Liber Genesis

1:1 in principio creauit Deus caelum et terram 1:2 terra 1:1 In the beginning God created the heaven and the earth. 1:2 And the earth was without form, and void; and darkness was upon the face of the deep. And the Spirit of God moved upon the face of the waters.

autem erat inanis et uacua et tenebrae super faciem abyssi et spiritus Dei ferebatur super aquas 1:3 dixitque Deus fiat 1:3 And God said, Let there be light: and there was light. lux et facta est lux 1:4 et uidit Deus lucem quod esset bona 1:4 And God saw the light, that it was good: and God

divided the light from the darkness.

et diuisit lucem ac tenebras 1:5 appellauitque lucem diem 1:5 And God called the light Day, and the darkness he

called Night. And the evening and the morning were the first day.

et tenebras noctem factumque est uespere et mane dies unus 1:6 dixit quoque Deus fiat firmamentum in medio aquarum 1:6 And God said, Let there be a firmament in the midst

of the waters, and let it divide the waters from the waters.

et diuidat aquas ab aquis 1:7 et fecit Deus firmamentum di- 1:7 And God made the firmament, and divided the waters

which were under the firmament from the waters which were above the firmament: and it was so.

uisitque aquas quae erant sub firmamento ab his quae erant super firmamentum et factum est ita 1:8 uocauitque Deus 1:8 And God called the firmament Heaven. And the

evening and the morning were the second day.

firmamentum caelum et factum est uespere et mane dies secundus 1:9 dixit uero Deus congregentur aquae quae sub 1:9 And God said, Let the waters under the heaven be

gathered together unto one place, and let the dry land appear: and it was so.

caelo sunt in locum unum et appareat arida factumque est ita 1:10 et uocauit Deus aridam terram congregationesque aqua- 1:10 And God called the dry land Earth; and the gathering

together of the waters called he Seas: and God saw that it was good.

rum appellauit maria et uidit Deus quod esset bonum 1:11

1:11 And God said, Let the earth bring forth grass, the herb yielding seed, and the fruit tree yielding fruit after his kind, whose seed is in itself, upon the earth: and it was so.

et ait germinet terra herbam uirentem et facientem semen et lignum pomiferum faciens fructum iuxta genus suum cuius semen in semet ipso sit super terram et factum est ita 1:12 et 1:12 And the earth brought forth grass, and herb yielding

seed after his kind, and the tree yielding fruit, whose seed was in itself, after his kind: and God saw that it was good.

protulit terra herbam uirentem et adferentem semen iuxta genus suum lignumque faciens fructum et habens unumquodque sementem secundum speciem suam et uidit Deus quod esset bonum 1:13 factumque est uespere et mane dies ter- 1:13 And the evening and the morning were the third day. tius 1:14 dixit autem Deus fiant luminaria in firmamento 1:14 And God said, Let there be lights in the firmament of

the heaven to divide the day from the night; and let them be for signs, and for seasons, and for days, and years:

caeli ut diuidant diem ac noctem et sint in signa et tempora et dies et annos 1:15 ut luceant in firmamento caeli et inlumi- 1:15 And let them be for lights in the firmament of the

heaven to give light upon the earth: and it was so.

nent terram et factum est ita 1:16 fecitque Deus duo magna 1:16 And God made two great lights; the greater light to

rule the day, and the lesser light to rule the night: he made the stars also.

luminaria luminare maius ut praeesset diei et luminare minus ut praeesset nocti et stellas 1:17 et posuit eas in firma- 1:17 And God set them in the firmament of the heaven to

give light upon the earth,

mento caeli ut lucerent super terram 1:18 et praeessent diei 1:18 And to rule over the day and over the night, and to

divide the light from the darkness: and God saw that it was good.

ac nocti et diuiderent lucem ac tenebras et uidit Deus quod esset bonum 1:19 et factum est uespere et mane dies quar- 1:19 And the evening and the morning were the fourth day. tus 1:20 dixit etiam Deus producant aquae reptile animae 1:20 And God said, Let the waters bring forth abundantly

the moving creature that hath life, and fowl that may fly above the earth in the open firmament of heaven.

uiuentis et uolatile super terram sub firmamento caeli 1:21

1:21 And God created great whales, and every living creature that moveth, which the waters brought forth abundantly, after their kind, and every winged fowl after his kind: and God saw that it was good.

Biblia Sacra Vulgata

creauitque Deus cete grandia et omnem animam uiuentem atque motabilem quam produxerant aquae in species suas et omne uolatile secundum genus suum et uidit Deus quod

1:22 And God blessed them, saying, Be fruitful, and mul- esset bonum 1:22 benedixitque eis dicens crescite et multi-

tiply, and fill the waters in the seas, and let fowl multiply in the earth.

plicamini et replete aquas maris auesque multiplicentur su-

1:23 And the evening and the morning were the fifth day. per terram 1:23 et factum est uespere et mane dies quintus 1:24 And God said, Let the earth bring forth the living 1:24 dixit quoque Deus producat terra animam uiuentem in

creature after his kind, cattle, and creeping thing, and beast of the earth after his kind: and it was so.

genere suo iumenta et reptilia et bestias terrae secundum

1:25 And God made the beast of the earth after his kind, species suas factumque est ita 1:25 et fecit Deus bestias

and cattle after their kind, and every thing that creepeth upon the earth after his kind: and God saw that it was

good. terrae iuxta species suas et iumenta et omne reptile terrae 1:26 And God said, Let us make man in our image, after in genere suo et uidit Deus quod esset bonum 1:26 et ait

our likeness: and let them have dominion over the fish of the sea, and over the fowl of the air, and over the cattle, and over all the earth, and over every creeping thing that creepeth upon the earth.

faciamus hominem ad imaginem et similitudinem nostram et praesit piscibus maris et uolatilibus caeli et bestiis uniuer-

1:27 So God created man in his own image, in the image saeque terrae omnique reptili quod mouetur in terra 1:27 et

of God created he him; male and female created he them.

creauit Deus hominem ad imaginem suam ad imaginem Dei

1:28 And God blessed them, and God said unto them, Be creauit illum masculum et feminam creauit eos 1:28 bene-

fruitful, and multiply, and replenish the earth, and subdue it: and have dominion over the fish of the sea, and over the fowl of the air, and over every living thing that moveth upon the earth.

dixitque illis Deus et ait crescite et multiplicamini et replete terram et subicite eam et dominamini piscibus maris et uolatilibus caeli et uniuersis animantibus quae mouentur super

1:29 And God said, Behold, I have given you every herb terram 1:29 dixitque Deus ecce dedi uobis omnem herbam

bearing seed, which is upon the face of all the earth, and every tree, in the which is the fruit of a tree yielding seed;

to you it shall be for meat. adferentem semen super terram et uniuersa ligna quae ha-

bent in semet ipsis sementem generis sui ut sint uobis in

1:30 And to every beast of the earth, and to every fowl escam 1:30 et cunctis animantibus terrae omnique uolucri

of the air, and to every thing that creepeth upon the earth, wherein there is life, I have given every green herb for

meat: and it was so. caeli et uniuersis quae mouentur in terra et in quibus est an-1:31 And God saw every thing that he had made, and, be- ima uiuens ut habeant ad uescendum et factum est ita 1:31

hold, it was very good. And the evening and the morning were the sixth day.

uiditque Deus cuncta quae fecit et erant ualde bona et factum est uespere et mane dies sextus

2:1 Thus the heavens and the earth were finished, and all 2:1 igitur perfecti sunt caeli et terra et omnis ornatus eo-

the host of them.

2:2 And on the seventh day God ended his work which he rum 2:2 conpleuitque Deus die septimo opus suum quod

had made; and he rested on the seventh day from all his work which he had made.

fecerat et requieuit die septimo ab uniuerso opere quod pat-

2:3 And God blessed the seventh day, and sanctified it: rarat 2:3 et benedixit diei septimo et sanctificauit illum quia

because that in it he had rested from all his work which God created and made.

in ipso cessauerat ab omni opere suo quod creauit Deus ut

2:4 These are the generations of the heavens and of the faceret 2:4 istae generationes caeli et terrae quando creatae

earth when they were created, in the day that the LORD God made the earth and the heavens,

Liber Genesis

sunt in die quo fecit Dominus Deus caelum et terram 2:5 2:5 And every plant of the field before it was in the earth, and every herb of the field before it grew: for the LORD God had not caused it to rain upon the earth, and there was

et omne uirgultum agri antequam oreretur in terra omnem- not a man to till the ground. que herbam regionis priusquam germinaret non enim pluerat Dominus Deus super terram et homo non erat qui operaretur terram 2:6 sed fons ascendebat e terra inrigans uniuersam 2:6 But there went up a mist from the earth, and watered

the whole face of the ground.

superficiem terrae 2:7 formauit igitur Dominus Deus ho- 2:7 And the LORD God formed man of the dust of the

ground, and breathed into his nostrils the breath of life; and man became a living soul.

minem de limo terrae et inspirauit in faciem eius spiraculum uitae et factus est homo in animam uiuentem 2:8 plantaue- 2:8 And the LORD God planted a garden eastward in

Eden; and there he put the man whom he had formed.

rat autem Dominus Deus paradisum uoluptatis a principio in quo posuit hominem quem formauerat 2:9 produxitque 2:9 And out of the ground made the LORD God to grow

every tree that is pleasant to the sight, and good for food; the tree of life also in the midst of the garden, and the tree

Dominus Deus de humo omne lignum pulchrum uisu et ad of knowledge of good and evil. uescendum suaue lignum etiam uitae in medio paradisi lignumque scientiae boni et mali 2:10 et fluuius egrediebatur 2:10 And a river went out of Eden to water the garden; and

from thence it was parted, and became into four heads.

de loco uoluptatis ad inrigandum paradisum qui inde diuiditur in quattuor capita 2:11 nomen uni Phison ipse est qui 2:11 The name of the first is Pison: that is it which com-

passeth the whole land of Havilah, where there is gold;

circuit omnem terram Euilat ubi nascitur aurum 2:12 et au- 2:12 And the gold of that land is good: there is bdellium

and the onyx stone.

rum terrae illius optimum est ibique inuenitur bdellium et lapis onychinus 2:13 et nomen fluuio secundo Geon ipse 2:13 And the name of the second river is Gihon: the same

is it that compasseth the whole land of Ethiopia.

est qui circuit omnem terram Aethiopiae 2:14 nomen uero 2:14 And the name of the third river is Hiddekel: that is

it which goeth toward the east of Assyria. And the fourth river is Euphrates.

fluminis tertii Tigris ipse uadit contra Assyrios fluuius autem quartus ipse est Eufrates 2:15 tulit ergo Dominus Deus 2:15 And the LORD God took the man, and put him into

the garden of Eden to dress it and to keep it.

hominem et posuit eum in paradiso uoluptatis ut operaretur et custodiret illum 2:16 praecepitque ei dicens ex omni 2:16 And the LORD God commanded the man, saying, Of

every tree of the garden thou mayest freely eat:

ligno paradisi comede 2:17 de ligno autem scientiae boni 2:17 But of the tree of the knowledge of good and evil,

thou shalt not eat of it: for in the day that thou eatest thereof thou shalt surely die.

et mali ne comedas in quocumque enim die comederis ex eo morte morieris 2:18 dixit quoque Dominus Deus non est 2:18 And the LORD God said, It is not good that the man

should be alone; I will make him an help meet for him.

bonum esse hominem solum faciamus ei adiutorium similem sui 2:19 formatis igitur Dominus Deus de humo cunc- 2:19 And out of the ground the LORD God formed every

beast of the field, and every fowl of the air; and brought them unto Adam to see what he would call them: and whatsoever Adam called every living creature, that was the name thereof.

tis animantibus terrae et uniuersis uolatilibus caeli adduxit ea ad Adam ut uideret quid uocaret ea omne enim quod uocauit Adam animae uiuentis ipsum est nomen eius 2:20 ap- 2:20 And Adam gave names to all cattle, and to the fowl of

the air, and to every beast of the field; but for Adam there was not found an help meet for him.

pellauitque Adam nominibus suis cuncta animantia et uniuersa uolatilia caeli et omnes bestias terrae Adam uero non

Biblia Sacra Vulgata

2:21 And the LORD God caused a deep sleep to fall upon inueniebatur adiutor similis eius 2:21 inmisit ergo Domi-

Adam, and he slept: and he took one of his ribs, and closed up the flesh instead thereof;

nus Deus soporem in Adam cumque obdormisset tulit unam

2:22 And the rib, which the LORD God had taken from de costis eius et repleuit carnem pro ea 2:22 et aedificauit

man, made he a woman, and brought her unto the man.

Dominus Deus costam quam tulerat de Adam in mulierem et

2:23 And Adam said, This is now bone of my bones, and adduxit eam ad Adam 2:23 dixitque Adam hoc nunc os ex

flesh of my flesh: she shall be called Woman, because she was taken out of Man.

ossibus meis et caro de carne mea haec uocabitur uirago quo-

2:24 Therefore shall a man leave his father and his mother, niam de uiro sumpta est 2:24 quam ob rem relinquet homo and shall cleave unto his wife: and they shall be one flesh.

patrem suum et matrem et adherebit uxori suae et erunt duo

2:25 And they were both naked, the man and his wife, and in carne una 2:25 erant autem uterque nudi Adam scilicet

were not ashamed.

et uxor eius et non erubescebant

3:1 Now the serpent was more subtil than any beast of the 3:1 sed et serpens erat callidior cunctis animantibus terrae

field which the LORD God had made. And he said unto the woman, Yea, hath God said, Ye shall not eat of every

tree of the garden? quae fecerat Dominus Deus qui dixit ad mulierem cur prae-

cepit uobis Deus ut non comederetis de omni ligno paradisi

3:2 And the woman said unto the serpent, We may eat of 3:2 cui respondit mulier de fructu lignorum quae sunt in pa-

the fruit of the trees of the garden:

3:3 But of the fruit of the tree which is in the midst of the radiso uescemur 3:3 de fructu uero ligni quod est in medio

garden, God hath said, Ye shall not eat of it, neither shall ye touch it, lest ye die.

paradisi praecepit nobis Deus ne comederemus et ne tange-

3:4 And the serpent said unto the woman, Ye shall not remus illud ne forte moriamur 3:4 dixit autem serpens ad

surely die: 3:5 For God doth know that in the day ye eat thereof, then mulierem nequaquam morte moriemini 3:5 scit enim Deus your eyes shall be opened, and ye shall be as gods, knowing good and evil.

quod in quocumque die comederitis ex eo aperientur oculi

3:6 And when the woman saw that the tree was good for uestri et eritis sicut dii scientes bonum et malum 3:6 uidit

food, and that it was pleasant to the eyes, and a tree to be desired to make one wise, she took of the fruit thereof, and did eat, and gave also unto her husband with her; and he did eat.

igitur mulier quod bonum esset lignum ad uescendum et pulchrum oculis aspectuque delectabile et tulit de fructu illius

3:7 And the eyes of them both were opened, and they knew et comedit deditque uiro suo qui comedit 3:7 et aperti sunt

that they were naked; and they sewed fig leaves together, and made themselves aprons.

oculi amborum cumque cognouissent esse se nudos consue-

3:8 And they heard the voice of the LORD God walking runt folia ficus et fecerunt sibi perizomata 3:8 et cum audis-

in the garden in the cool of the day: and Adam and his wife hid themselves from the presence of the LORD God

amongst the trees of the garden. sent uocem Domini Dei deambulantis in paradiso ad auram

post meridiem abscondit se Adam et uxor eius a facie Do-

3:9 And the LORD God called unto Adam, and said unto mini Dei in medio ligni paradisi 3:9 uocauitque Dominus

him, Where art thou? 3:10 And he said, I heard thy voice in the garden, and I Deus Adam et dixit ei ubi es 3:10 qui ait uocem tuam auwas afraid, because I was naked; and I hid myself.

diui in paradiso et timui eo quod nudus essem et abscondi

3:11 And he said, Who told thee that thou wast naked? me 3:11 cui dixit quis enim indicauit tibi quod nudus es-

Hast thou eaten of the tree, whereof I commanded thee that thou shouldest not eat?

ses nisi quod ex ligno de quo tibi praeceperam ne comederes

3:12 And the man said, The woman whom thou gavest to comedisti 3:12 dixitque Adam mulier quam dedisti sociam

be with me, she gave me of the tree, and I did eat.

Liber Genesis

mihi dedit mihi de ligno et comedi 3:13 et dixit Dominus 3:13 And the LORD God said unto the woman, What is

this that thou hast done? And the woman said, The serpent beguiled me, and I did eat.

Deus ad mulierem quare hoc fecisti quae respondit serpens decepit me et comedi 3:14 et ait Dominus Deus ad serpen- 3:14 And the LORD God said unto the serpent, Because

thou hast done this, thou art cursed above all cattle, and above every beast of the field; upon thy belly shalt thou

tem quia fecisti hoc maledictus es inter omnia animantia et go, and dust shalt thou eat all the days of thy life: bestias terrae super pectus tuum gradieris et terram comedes cunctis diebus uitae tuae 3:15 inimicitias ponam inter te et 3:15 And I will put enmity between thee and the woman,

and between thy seed and her seed; it shall bruise thy head, and thou shalt bruise his heel.

mulierem et semen tuum et semen illius ipsa conteret caput tuum et tu insidiaberis calcaneo eius 3:16 mulieri quoque 3:16 Unto the woman he said, I will greatly multiply thy

sorrow and thy conception; in sorrow thou shalt bring forth children; and thy desire shall be to thy husband, and he

dixit multiplicabo aerumnas tuas et conceptus tuos in do- shall rule over thee. lore paries filios et sub uiri potestate eris et ipse dominabitur tui 3:17 ad Adam uero dixit quia audisti uocem uxoris 3:17 And unto Adam he said, Because thou hast hearkened

unto the voice of thy wife, and hast eaten of the tree, of which I commanded thee, saying, Thou shalt not eat of it: cursed is the ground for thy sake; in sorrow shalt thou eat of it all the days of thy life;

tuae et comedisti de ligno ex quo praeceperam tibi ne comederes maledicta terra in opere tuo in laboribus comedes eam cunctis diebus uitae tuae 3:18 spinas et tribulos ger- 3:18 Thorns also and thistles shall it bring forth to thee;

and thou shalt eat the herb of the field;

minabit tibi et comedes herbas terrae 3:19 in sudore uultus 3:19 In the sweat of thy face shalt thou eat bread, till thou

return unto the ground; for out of it wast thou taken: for dust thou art, and unto dust shalt thou return.

tui uesceris pane donec reuertaris in terram de qua sumptus es quia puluis es et in puluerem reuerteris 3:20 et uocauit 3:20 And Adam called his wife’s name Eve; because she

was the mother of all living.

Adam nomen uxoris suae Haua eo quod mater esset cunctorum uiuentium 3:21 fecit quoque Dominus Deus Adam 3:21 Unto Adam also and to his wife did the LORD God

make coats of skins, and clothed them.

et uxori eius tunicas pellicias et induit eos 3:22 et ait ecce 3:22 And the LORD God said, Behold, the man is become

as one of us, to know good and evil: and now, lest he put forth his hand, and take also of the tree of life, and eat, and

Adam factus est quasi unus ex nobis sciens bonum et malum live for ever: nunc ergo ne forte mittat manum suam et sumat etiam de ligno uitae et comedat et uiuat in aeternum 3:23 emisit eum 3:23 Therefore the LORD God sent him forth from the

garden of Eden, to till the ground from whence he was taken.

Dominus Deus de paradiso uoluptatis ut operaretur terram de qua sumptus est 3:24 eiecitque Adam et conlocauit ante 3:24 So he drove out the man; and he placed at the east of

the garden of Eden Cherubims, and a flaming sword which turned every way, to keep the way of the tree of life.

paradisum uoluptatis cherubin et flammeum gladium atque uersatilem ad custodiendam uiam ligni uitae

4:1 Adam uero cognouit Hauam uxorem suam quae con- 4:1 And Adam knew Eve his wife; and she conceived, and

bare Cain, and said, I have gotten a man from the LORD.

cepit et peperit Cain dicens possedi hominem per Dominum 4:2 rursusque peperit fratrem eius Abel fuit autem Abel pas- 4:2 And she again bare his brother Abel. And Abel was a

keeper of sheep, but Cain was a tiller of the ground.

tor ouium et Cain agricola 4:3 factum est autem post mul- 4:3 And in process of time it came to pass, that Cain

brought of the fruit of the ground an offering unto the LORD.

tos dies ut offerret Cain de fructibus terrae munera Domino 4:4 Abel quoque obtulit de primogenitis gregis sui et de ad- 4:4 And Abel, he also brought of the firstlings of his flock

and of the fat thereof. And the LORD had respect unto Abel and to his offering:

Biblia Sacra Vulgata

ipibus eorum et respexit Dominus ad Abel et ad munera eius

4:5 But unto Cain and to his offering he had not respect. 4:5 ad Cain uero et ad munera illius non respexit iratusque est

And Cain was very wroth, and his countenance fell.

4:6 And the LORD said unto Cain, Why art thou wroth? Cain uehementer et concidit uultus eius 4:6 dixitque Domi-

and why is thy countenance fallen? 4:7 If thou doest well, shalt thou not be accepted? and if nus ad eum quare maestus es et cur concidit facies tua 4:7 thou doest not well, sin lieth at the door. And unto thee shall be his desire, and thou shalt rule over him.

nonne si bene egeris recipies sin autem male statim in foribus peccatum aderit sed sub te erit appetitus eius et tu domi-

4:8 And Cain talked with Abel his brother: and it came to naberis illius 4:8 dixitque Cain ad Abel fratrem suum egre-

pass, when they were in the field, that Cain rose up against Abel his brother, and slew him.

diamur foras cumque essent in agro consurrexit Cain aduer-

4:9 And the LORD said unto Cain, Where is Abel thy sus Abel fratrem suum et interfecit eum 4:9 et ait Dominus

brother? And he said, I know not: Am I my brother’s keeper?

ad Cain ubi est Abel frater tuus qui respondit nescio num cu-

4:10 And he said, What hast thou done? the voice of thy stos fratris mei sum 4:10 dixitque ad eum quid fecisti uox

brother’s blood crieth unto me from the ground.

4:11 And now art thou cursed from the earth, which hath sanguinis fratris tui clamat ad me de terra 4:11 nunc igitur

opened her mouth to receive thy brother’s blood from thy hand;

maledictus eris super terram quae aperuit os suum et suscepit

4:12 When thou tillest the ground, it shall not henceforth sanguinem fratris tui de manu tua 4:12 cum operatus fueris yield unto thee her strength; a fugitive and a vagabond shalt thou be in the earth.

eam non dabit tibi fructus suos uagus et profugus eris super

4:13 And Cain said unto the LORD, My punishment is terram 4:13 dixitque Cain ad Dominum maior est iniquitas

greater than I can bear.

4:14 Behold, thou hast driven me out this day from the mea quam ut ueniam merear 4:14 ecce eicis me hodie a fa-

face of the earth; and from thy face shall I be hid; and I shall be a fugitive and a vagabond in the earth; and it shall

come to pass, that every one that findeth me shall slay me. cie terrae et a facie tua abscondar et ero uagus et profugus in 4:15 And the LORD said unto him, Therefore whosoever terra omnis igitur qui inuenerit me occidet me 4:15 dixit-

slayeth Cain, vengeance shall be taken on him sevenfold. And the LORD set a mark upon Cain, lest any finding him

should kill him. que ei Dominus nequaquam ita fiet sed omnis qui occiderit

Cain septuplum punietur posuitque Dominus Cain signum

4:16 And Cain went out from the presence of the LORD, ut non eum interficeret omnis qui inuenisset eum 4:16 eg-

and dwelt in the land of Nod, on the east of Eden.

ressusque Cain a facie Domini habitauit in terra profugus ad

4:17 And Cain knew his wife; and she conceived, and bare orientalem plagam Eden 4:17 cognouit autem Cain uxorem

Enoch: and he builded a city, and called the name of the city, after the name of his son, Enoch.

suam quae concepit et peperit Enoch et aedificauit ciuita-

4:18 And unto Enoch was born Irad: and Irad begat Mehu- tem uocauitque nomen eius ex nomine filii sui Enoch 4:18

jael: and Mehujael begat Methusael: and Methusael begat Lamech.

porro Enoch genuit Irad et Irad genuit Mauiahel et Mauia-

4:19 And Lamech took unto him two wives: the name of hel genuit Matusahel et Matusahel genuit Lamech 4:19 qui

the one was Adah, and the name of the other Zillah.

accepit uxores duas nomen uni Ada et nomen alteri Sella

4:20 And Adah bare Jabal: he was the father of such as 4:20 genuitque Ada Iabel qui fuit pater habitantium in ten-

dwell in tents, and of such as have cattle.

4:21 And his brother’s name was Jubal: he was the father toriis atque pastorum 4:21 et nomen fratris eius Iubal ipse

of all such as handle the harp and organ. 4:22 And Zillah, she also bare Tubalcain, an instructer of fuit pater canentium cithara et organo 4:22 Sella quoque every artificer in brass and iron: and the sister of Tubalcain was Naamah.

genuit Thubalcain qui fuit malleator et faber in cuncta opera

Liber Genesis

aeris et ferri soror uero Thubalcain Noemma 4:23 dixitque 4:23 And Lamech said unto his wives, Adah and Zillah,

Hear my voice; ye wives of Lamech, hearken unto my speech: for I have slain a man to my wounding, and a

Lamech uxoribus suis Adae et Sellae audite uocem meam young man to my hurt. uxores Lamech auscultate sermonem meum quoniam occidi uirum in uulnus meum et adulescentulum in liuorem meum

4:24 septuplum ultio dabitur de Cain de Lamech uero sep- 4:24 If Cain shall be avenged sevenfold, truly Lamech sev-

enty and sevenfold.

tuagies septies 4:25 cognouit quoque adhuc Adam uxorem 4:25 And Adam knew his wife again; and she bare a son,

and called his name Seth: For God, said she, hath appointed me another seed instead of Abel, whom Cain slew.

suam et peperit filium uocauitque nomen eius Seth dicens posuit mihi Deus semen aliud pro Abel quem occidit Cain 4:26 sed et Seth natus est filius quem uocauit Enos iste coepit 4:26 And to Seth, to him also there was born a son; and

he called his name Enos: then began men to call upon the name of the LORD.

inuocare nomen Domini

5:1 hic est liber generationis Adam in die qua creauit Deus 5:1 This is the book of the generations of Adam. In the

day that God created man, in the likeness of God made he him;

hominem ad similitudinem Dei fecit illum 5:2 masculum

5:2 Male and female created he them; and blessed them, and called their name Adam, in the day when they were created.

et feminam creauit eos et benedixit illis et uocauit nomen

eorum Adam in die qua creati sunt 5:3 uixit autem Adam 5:3 And Adam lived an hundred and thirty years, and begat a son in his own likeness, and after his image; and called his name Seth:

centum triginta annis et genuit ad similitudinem et imaginem suam uocauitque nomen eius Seth 5:4 et facti sunt dies Ad- 5:4 And the days of Adam after he had begotten Seth were

eight hundred years: and he begat sons and daughters:

am postquam genuit Seth octingenti anni genuitque filios et filias 5:5 et factum est omne tempus quod uixit Adam anni 5:5 And all the days that Adam lived were nine hundred

and thirty years: and he died.

nongenti triginta et mortuus est 5:6 uixit quoque Seth cen- 5:6 And Seth lived an hundred and five years, and begat

Enos:

tum quinque annos et genuit Enos 5:7 uixitque Seth post- 5:7 And Seth lived after he begat Enos eight hundred and

seven years, and begat sons and daughters:

quam genuit Enos octingentis septem annis genuitque filios et filias 5:8 et facti sunt omnes dies Seth nongentorum duo- 5:8 And all the days of Seth were nine hundred and twelve

years: and he died.

decim annorum et mortuus est 5:9 uixit uero Enos nona- 5:9 And Enos lived ninety years, and begat Cainan: ginta annis et genuit Cainan 5:10 post cuius ortum uixit 5:10 And Enos lived after he begat Cainan eight hundred

and fifteen years, and begat sons and daughters:

octingentis quindecim annis et genuit filios et filias 5:11 5:11 And all the days of Enos were nine hundred and five

years: and he died.

factique sunt omnes dies Enos nongentorum quinque annorum et mortuus est 5:12 uixit quoque Cainan septuaginta 5:12 And Cainan lived seventy years and begat Maha-

laleel:

annis et genuit Malalehel 5:13 et uixit Cainan postquam 5:13 And Cainan lived after he begat Mahalaleel eight

hundred and forty years, and begat sons and daughters:

genuit Malalehel octingentos quadraginta annos genuitque filios et filias 5:14 et facti sunt omnes dies Cainan nongenti 5:14 And all the days of Cainan were nine hundred and

ten years: and he died.

decem anni et mortuus est 5:15 uixit autem Malalehel sexa- 5:15 And Mahalaleel lived sixty and five years, and begat

Jared:

ginta quinque annos et genuit Iared 5:16 et uixit Malalehel 5:16 And Mahalaleel lived after he begat Jared eight hun-

dred and thirty years, and begat sons and daughters:

Biblia Sacra Vulgata

postquam genuit Iared octingentis triginta annis et genuit fi-5:17 And all the days of Mahalaleel were eight hundred lios et filias 5:17 et facti sunt omnes dies Malalehel octin-

ninety and five years: and he died.

5:18 And Jared lived an hundred sixty and two years, and genti nonaginta quinque anni et mortuus est 5:18 uixitque

he begat Enoch:

5:19 And Jared lived after he begat Enoch eight hundred Iared centum sexaginta duobus annis et genuit Enoch 5:19

years, and begat sons and daughters:

et uixit Iared postquam genuit Enoch octingentos annos et

5:20 And all the days of Jared were nine hundred sixty and genuit filios et filias 5:20 et facti sunt omnes dies Iared non-

two years: and he died.

5:21 And Enoch lived sixty and five years, and begat genti sexaginta duo anni et mortuus est 5:21 porro Enoch

Methuselah:

5:22 And Enoch walked with God after he begat Methuse- uixit sexaginta quinque annis et genuit Mathusalam 5:22

lah three hundred years, and begat sons and daughters:

et ambulauit Enoch cum Deo postquam genuit Mathusalam

5:23 And all the days of Enoch were three hundred sixty trecentis annis et genuit filios et filias 5:23 et facti sunt om-

and five years:

5:24 And Enoch walked with God: and he was not; for nes dies Enoch trecenti sexaginta quinque anni 5:24 ambu-

God took him.

5:25 And Methuselah lived an hundred eighty and seven lauitque cum Deo et non apparuit quia tulit eum Deus 5:25

years, and begat Lamech.

uixit quoque Mathusalam centum octoginta septem annos et

5:26 And Methuselah lived after he begat Lamech seven genuit Lamech 5:26 et uixit Mathusalam postquam genuit

hundred eighty and two years, and begat sons and daughters:

Lamech septingentos octoginta duos annos et genuit filios

5:27 And all the days of Methuselah were nine hundred et filias 5:27 et facti sunt omnes dies Mathusalae nongenti

sixty and nine years: and he died.

5:28 And Lamech lived an hundred eighty and two years, sexaginta nouem anni et mortuus est 5:28 uixit autem La-

and begat a son: 5:29 And he called his name Noah, saying, This same shall mech centum octoginta duobus annis et genuit filium 5:29 comfort us concerning our work and toil of our hands, because of the ground which the LORD hath cursed.

uocauitque nomen eius Noe dicens iste consolabitur nos ab operibus et laboribus manuum nostrarum in terra cui male-

5:30 And Lamech lived after he begat Noah five hundred dixit Dominus 5:30 uixitque Lamech postquam genuit Noe

ninety and five years, and begat sons and daughters:

quingentos nonaginta quinque annos et genuit filios et filias

5:31 And all the days of Lamech were seven hundred sev- 5:31 et facti sunt omnes dies Lamech septingenti septuaginta

enty and seven years: and he died.

5:32 And Noah was five hundred years old: and Noah be- septem anni et mortuus est 5:32 Noe uero cum quingento-

gat Shem, Ham, and Japheth.

rum esset annorum genuit Sem et Ham et Iafeth

6:1 And it came to pass, when men began to multiply on 6:1 cumque coepissent homines multiplicari super terram

the face of the earth, and daughters were born unto them,

6:2 That the sons of God saw the daughters of men that et filias procreassent 6:2 uidentes filii Dei filias eorum quod

they were fair; and they took them wives of all which they chose.

essent pulchrae acceperunt uxores sibi ex omnibus quas ele-

6:3 And the LORD said, My spirit shall not always strive gerant 6:3 dixitque Deus non permanebit spiritus meus in

with man, for that he also is flesh: yet his days shall be an hundred and twenty years.

homine in aeternum quia caro est eruntque dies illius cen-

6:4 There were giants in the earth in those days; and also tum uiginti annorum 6:4 gigantes autem erant super terram

after that, when the sons of God came in unto the daughters of men, and they bare children to them, the same be-

came mighty men which were of old, men of renown. in diebus illis postquam enim ingressi sunt filii Dei ad filias hominum illaeque genuerunt isti sunt potentes a saeculo uiri

6:5 And God saw that the wickedness of man was great in famosi 6:5 uidens autem Deus quod multa malitia homi-

the earth, and that every imagination of the thoughts of his heart was only evil continually.

Liber Genesis

num esset in terra et cuncta cogitatio cordis intenta esset ad malum omni tempore 6:6 paenituit eum quod hominem fe- 6:6 And it repented the LORD that he had made man on

the earth, and it grieved him at his heart.

cisset in terra et tactus dolore cordis intrinsecus 6:7 delebo 6:7 And the LORD said, I will destroy man whom I have

created from the face of the earth; both man, and beast, and the creeping thing, and the fowls of the air; for it repenteth

inquit hominem quem creaui a facie terrae ab homine usque me that I have made them. ad animantia a reptili usque ad uolucres caeli paenitet enim me fecisse eos 6:8 Noe uero inuenit gratiam coram Domino 6:8 But Noah found grace in the eyes of the LORD. 6:9 hae generationes Noe Noe uir iustus atque perfectus fuit 6:9 These are the generations of Noah: Noah was a just

man and perfect in his generations, and Noah walked with God.

in generationibus suis cum Deo ambulauit 6:10 et genuit

6:10 And Noah begat three sons, Shem, Ham, and Japheth.

tres filios Sem Ham et Iafeth 6:11 corrupta est autem terra 6:11 The earth also was corrupt before God, and the earth

was filled with violence.

coram Deo et repleta est iniquitate 6:12 cumque uidisset 6:12 And God looked upon the earth, and, behold, it was

corrupt; for all flesh had corrupted his way upon the earth.

Deus terram esse corruptam omnis quippe caro corruperat uiam suam super terram 6:13 dixit ad Noe finis uniuersae 6:13 And God said unto Noah, The end of all flesh is come

before me; for the earth is filled with violence through them; and, behold, I will destroy them with the earth.

carnis uenit coram me repleta est terra iniquitate a facie eorum et ego disperdam eos cum terra 6:14 fac tibi arcam de 6:14 Make thee an ark of gopher wood; rooms shalt thou

make in the ark, and shalt pitch it within and without with pitch.

lignis leuigatis mansiunculas in arca facies et bitumine linies intrinsecus et extrinsecus 6:15 et sic facies eam trecento- 6:15 And this is the fashion which thou shalt make it of:

The length of the ark shall be three hundred cubits, the breadth of it fifty cubits, and the height of it thirty cubits.

rum cubitorum erit longitudo arcae quinquaginta cubitorum latitudo et triginta cubitorum altitudo illius 6:16 fenestram 6:16 A window shalt thou make to the ark, and in a cubit

shalt thou finish it above; and the door of the ark shalt thou set in the side thereof; with lower, second, and third stories

in arca facies et in cubito consummabis summitatem ostium shalt thou make it. autem arcae pones ex latere deorsum cenacula et tristega facies in ea 6:17 ecce ego adducam diluuii aquas super terram 6:17 And, behold, I, even I, do bring a flood of waters

upon the earth, to destroy all flesh, wherein is the breath of life, from under heaven; and every thing that is in the

ut interficiam omnem carnem in qua spiritus uitae est subter earth shall die. caelum uniuersa quae in terra sunt consumentur 6:18 po- 6:18 But with thee will I establish my covenant; and thou

shalt come into the ark, thou, and thy sons, and thy wife, and thy sons’ wives with thee.

namque foedus meum tecum et ingredieris arcam tu et filii tui uxor tua et uxores filiorum tuorum tecum 6:19 et ex 6:19 And of every living thing of all flesh, two of every

sort shalt thou bring into the ark, to keep them alive with thee; they shall be male and female.

cunctis animantibus uniuersae carnis bina induces in arcam

ut uiuant tecum masculini sexus et feminini 6:20 de uolu- 6:20 Of fowls after their kind, and of cattle after their kind, of every creeping thing of the earth after his kind, two of every sort shall come unto thee, to keep them alive.

cribus iuxta genus suum et de iumentis in genere suo et ex omni reptili terrae secundum genus suum bina de omnibus ingredientur tecum ut possint uiuere 6:21 tolles igitur te- 6:21 And take thou unto thee of all food that is eaten, and

thou shalt gather it to thee; and it shall be for food for thee, and for them.

cum ex omnibus escis quae mandi possunt et conportabis apud te et erunt tam tibi quam illis in cibum 6:22 fecit ergo 6:22 Thus did Noah; according to all that God commanded

him, so did he.

Noe omnia quae praeceperat illi Deus

Biblia Sacra Vulgata

7:1 And the LORD said unto Noah, Come thou and all thy 7:1 dixitque Dominus ad eum ingredere tu et omnis domus

house into the ark; for thee have I seen righteous before me in this generation.

tua arcam te enim uidi iustum coram me in generatione hac

7:2 Of every clean beast thou shalt take to thee by sevens, 7:2 ex omnibus animantibus mundis tolle septena septena the male and his female: and of beasts that are not clean by two, the male and his female.

masculum et feminam de animantibus uero non mundis duo

7:3 Of fowls also of the air by sevens, the male and the duo masculum et feminam 7:3 sed et de uolatilibus caeli se-

female; to keep seed alive upon the face of all the earth.

ptena septena masculum et feminam ut saluetur semen super

7:4 For yet seven days, and I will cause it to rain upon faciem uniuersae terrae 7:4 adhuc enim et post dies septem

the earth forty days and forty nights; and every living substance that I have made will I destroy from off the face of

the earth. ego pluam super terram quadraginta diebus et quadraginta

noctibus et delebo omnem substantiam quam feci de super-

7:5 And Noah did according unto all that the LORD com- ficie terrae 7:5 fecit ergo Noe omnia quae mandauerat ei

manded him.

7:6 And Noah was six hundred years old when the flood Dominus 7:6 eratque sescentorum annorum quando diluuii

of waters was upon the earth.

7:7 And Noah went in, and his sons, and his wife, and his aquae inundauerunt super terram 7:7 et ingressus est Noe

sons’ wives with him, into the ark, because of the waters of the flood.

et filii eius uxor eius et uxores filiorum eius cum eo in arcam

7:8 Of clean beasts, and of beasts that are not clean, and propter aquas diluuii 7:8 de animantibus quoque mundis et

of fowls, and of every thing that creepeth upon the earth,

inmundis et de uolucribus et ex omni quod mouetur super

7:9 There went in two and two unto Noah into the ark, the terram 7:9 duo et duo ingressa sunt ad Noe in arcam mamale and the female, as God had commanded Noah. 7:10 And it came to pass after seven days, that the waters sculus et femina sicut praeceperat Deus Noe 7:10 cumque of the flood were upon the earth.

transissent septem dies aquae diluuii inundauerunt super ter-

7:11 In the six hundredth year of Noah’s life, in the second ram 7:11 anno sescentesimo uitae Noe mense secundo semonth, the seventeenth day of the month, the same day were all the fountains of the great deep broken up, and the

windows of heaven were opened. ptimodecimo die mensis rupti sunt omnes fontes abyssi ma-7:12 And the rain was upon the earth forty days and forty gnae et cataractae caeli apertae sunt 7:12 et facta est plu-

nights.

uia super terram quadraginta diebus et quadraginta noctibus

7:13 In the selfsame day entered Noah, and Shem, and 7:13 in articulo diei illius ingressus est Noe et Sem et Ham et

Ham, and Japheth, the sons of Noah, and Noah’s wife, and the three wives of his sons with them, into the ark;

Iafeth filii eius uxor illius et tres uxores filiorum eius cum eis

7:14 They, and every beast after his kind, and all the cat- in arcam 7:14 ipsi et omne animal secundum genus suum tle after their kind, and every creeping thing that creepeth upon the earth after his kind, and every fowl after his kind,

every bird of every sort. uniuersaque iumenta in genus suum et omne quod moue-

tur super terram in genere suo cunctumque uolatile secun-

7:15 And they went in unto Noah into the ark, two and two dum genus suum uniuersae aues omnesque uolucres 7:15 of all flesh, wherein is the breath of life.

ingressae sunt ad Noe in arcam bina et bina ex omni carne

7:16 And they that went in, went in male and female of all in qua erat spiritus uitae 7:16 et quae ingressa sunt mas-

flesh, as God had commanded him: and the LORD shut him in.

culus et femina ex omni carne introierunt sicut praeceperat

7:17 And the flood was forty days upon the earth; and the ei Deus et inclusit eum Dominus de foris 7:17 factumque

waters increased, and bare up the ark, and it was lift up above the earth.

est diluuium quadraginta diebus super terram et multiplica-

Liber Genesis

tae sunt aquae et eleuauerunt arcam in sublime a terra 7:18 7:18 And the waters prevailed, and were increased greatly

upon the earth; and the ark went upon the face of the waters.

uehementer inundauerunt et omnia repleuerunt in superficie terrae porro arca ferebatur super aquas 7:19 et aquae prae- 7:19 And the waters prevailed exceedingly upon the earth;

and all the high hills, that were under the whole heaven, were covered.

ualuerunt nimis super terram opertique sunt omnes montes excelsi sub uniuerso caelo 7:20 quindecim cubitis altior fuit 7:20 Fifteen cubits upward did the waters prevail; and the

mountains were covered.

aqua super montes quos operuerat 7:21 consumptaque est 7:21 And all flesh died that moved upon the earth, both

of fowl, and of cattle, and of beast, and of every creeping thing that creepeth upon the earth, and every man:

omnis caro quae mouebatur super terram uolucrum animantium bestiarum omniumque reptilium quae reptant super ter-

ram uniuersi homines 7:22 et cuncta in quibus spiraculum 7:22 All in whose nostrils was the breath of life, of all that

was in the dry land, died.

uitae est in terra mortua sunt 7:23 et deleuit omnem sub- 7:23 And every living substance was destroyed which was

upon the face of the ground, both man, and cattle, and the creeping things, and the fowl of the heaven; and they were destroyed from the earth: and Noah only remained alive, and they that were with him in the ark.

stantiam quae erat super terram ab homine usque ad pecus tam reptile quam uolucres caeli et deleta sunt de terra re-

mansit autem solus Noe et qui cum eo erant in arca 7:24 7:24 And the waters prevailed upon the earth an hundred

and fifty days.

obtinueruntque aquae terras centum quinquaginta diebus

8:1 recordatus autem Deus Noe cunctarumque animan- 8:1 And God remembered Noah, and every living thing,

and all the cattle that was with him in the ark: and God made a wind to pass over the earth, and the waters ass-

tium et omnium iumentorum quae erant cum eo in arca ad- waged; duxit spiritum super terram et inminutae sunt aquae 8:2 8:2 The fountains also of the deep and the windows of

heaven were stopped, and the rain from heaven was restrained;

et clausi sunt fontes abyssi et cataractae caeli et prohibitae sunt pluuiae de caelo 8:3 reuersaeque aquae de terra euntes 8:3 And the waters returned from off the earth continually:

and after the end of the hundred and fifty days the waters were abated.

et redeentes et coeperunt minui post centum quinquaginta dies 8:4 requieuitque arca mense septimo uicesima sep- 8:4 And the ark rested in the seventh month, on the seven-

teenth day of the month, upon the mountains of Ararat.

tima die mensis super montes Armeniae 8:5 at uero aquae 8:5 And the waters decreased continually until the tenth month: in the tenth month, on the first day of the month, were the tops of the mountains seen.

ibant et decrescebant usque ad decimum mensem decimo enim mense prima die mensis apparuerunt cacumina montium 8:6 cumque transissent quadraginta dies aperiens Noe 8:6 And it came to pass at the end of forty days, that Noah

opened the window of the ark which he had made:

fenestram arcae quam fecerat dimisit coruum 8:7 qui egre- 8:7 And he sent forth a raven, which went forth to and fro,

until the waters were dried up from off the earth.

diebatur et reuertebatur donec siccarentur aquae super ter-

ram 8:8 emisit quoque columbam post eum ut uideret si 8:8 Also he sent forth a dove from him, to see if the waters

were abated from off the face of the ground;

iam cessassent aquae super faciem terrae 8:9 quae cum non 8:9 But the dove found no rest for the sole of her foot, and

she returned unto him into the ark, for the waters were on the face of the whole earth: then he put forth his hand, and

inuenisset ubi requiesceret pes eius reuersa est ad eum in ar- took her, and pulled her in unto him into the ark. cam aquae enim erant super uniuersam terram extenditque manum et adprehensam intulit in arcam 8:10 expectatis au- 8:10 And he stayed yet other seven days; and again he sent

forth the dove out of the ark;

tem ultra septem diebus aliis rursum dimisit columbam ex

Biblia Sacra Vulgata

8:11 And the dove came in to him in the evening; and, lo, arca 8:11 at illa uenit ad eum ad uesperam portans ramum

in her mouth was an olive leaf pluckt off: so Noah knew that the waters were abated from off the earth.

oliuae uirentibus foliis in ore suo intellexit ergo Noe quod

8:12 And he stayed yet other seven days; and sent forth cessassent aquae super terram 8:12 expectauitque nihilo-

the dove; which returned not again unto him any more.

minus septem alios dies et emisit columbam quae non est

8:13 And it came to pass in the six hundredth and first year, reuersa ultra ad eum 8:13 igitur sescentesimo primo anno in the first month, the first day of the month, the waters were dried up from off the earth: and Noah removed the covering of the ark, and looked, and, behold, the face of the ground was dry.

primo mense prima die mensis inminutae sunt aquae super terram et aperiens Noe tectum arcae aspexit uiditque quod

8:14 And in the second month, on the seven and twentieth exsiccata esset superficies terrae 8:14 mense secundo se-

day of the month, was the earth dried.

8:15 And God spake unto Noah, saying, ptima et uicesima die mensis arefacta est terra 8:15 locu-

8:16 Go forth of the ark, thou, and thy wife, and thy sons, tus est autem Deus ad Noe dicens 8:16 egredere de arca tu and thy sons’ wives with thee. 8:17 Bring forth with thee every living thing that is with et uxor tua filii tui et uxores filiorum tuorum tecum 8:17 thee, of all flesh, both of fowl, and of cattle, and of every creeping thing that creepeth upon the earth; that they may breed abundantly in the earth, and be fruitful, and multiply upon the earth.

cuncta animantia quae sunt apud te ex omni carne tam in uolatilibus quam in bestiis et in uniuersis reptilibus quae reptant super terram educ tecum et ingredimini super ter-

8:18 And Noah went forth, and his sons, and his wife, and ram crescite et multiplicamini super eam 8:18 egressus est

his sons’ wives with him:

ergo Noe et filii eius uxor illius et uxores filiorum eius cum

8:19 Every beast, every creeping thing, and every fowl, eo 8:19 sed et omnia animantia iumenta et reptilia quae re-

and whatsoever creepeth upon the earth, after their kinds, went forth out of the ark.

punt super terram secundum genus suum arcam egressa sunt

8:20 And Noah builded an altar unto the LORD; and took 8:20 aedificauit autem Noe altare Domino et tollens de cunc-

of every clean beast, and of every clean fowl, and offered burnt offerings on the altar.

tis pecoribus et uolucribus mundis obtulit holocausta super

8:21 And the LORD smelled a sweet savour; and the altare 8:21 odoratusque est Dominus odorem suauitatis et

LORD said in his heart, I will not again curse the ground any more for man’s sake; for the imagination of man’s heart is evil from his youth; neither will I again smite any more every thing living, as I have done.

ait ad eum nequaquam ultra maledicam terrae propter homines sensus enim et cogitatio humani cordis in malum prona sunt ab adulescentia sua non igitur ultra percutiam omnem

8:22 While the earth remaineth, seedtime and harvest, and animantem sicut feci 8:22 cunctis diebus terrae sementis et

cold and heat, and summer and winter, and day and night shall not cease.

messis frigus et aestus aestas et hiemps nox et dies non requiescent

9:1 And God blessed Noah and his sons, and said unto 9:1 benedixitque Deus Noe et filiis eius et dixit ad eos cre-

them, Be fruitful, and multiply, and replenish the earth.

9:2 And the fear of you and the dread of you shall be upon scite et multiplicamini et implete terram 9:2 et terror uester

every beast of the earth, and upon every fowl of the air, upon all that moveth upon the earth, and upon all the fishes

of the sea; into your hand are they delivered. ac tremor sit super cuncta animalia terrae et super omnes

uolucres caeli cum uniuersis quae mouentur in terra omnes

9:3 Every moving thing that liveth shall be meat for you; pisces maris manui uestrae traditi sunt 9:3 et omne quod even as the green herb have I given you all things.

mouetur et uiuit erit uobis in cibum quasi holera uirentia tra-

Liber Genesis

didi uobis omnia 9:4 excepto quod carnem cum sanguine 9:4 But flesh with the life thereof, which is the blood

thereof, shall ye not eat.

non comedetis 9:5 sanguinem enim animarum uestrarum 9:5 And surely your blood of your lives will I require; at the hand of every beast will I require it, and at the hand of man; at the hand of every man’s brother will I require the

requiram de manu cunctarum bestiarum et de manu hominis life of man. de manu uiri et fratris eius requiram animam hominis 9:6 9:6 Whoso sheddeth man’s blood, by man shall his blood

be shed: for in the image of God made he man.

quicumque effuderit humanum sanguinem fundetur sanguis

illius ad imaginem quippe Dei factus est homo 9:7 uos au- 9:7 And you, be ye fruitful, and multiply; bring forth abun-

dantly in the earth, and multiply therein.

tem crescite et multiplicamini et ingredimini super terram et implete eam 9:8 haec quoque dixit Deus ad Noe et ad fi- 9:8 And God spake unto Noah, and to his sons with him,

saying,

lios eius cum eo 9:9 ecce ego statuam pactum meum uobi- 9:9 And I, behold, I establish my covenant with you, and

with your seed after you;

scum et cum semine uestro post uos 9:10 et ad omnem ani- 9:10 And with every living creature that is with you, of the fowl, of the cattle, and of every beast of the earth with you; from all that go out of the ark, to every beast of the

mam uiuentem quae est uobiscum tam in uolucribus quam earth. in iumentis et pecudibus terrae cunctis quae egressa sunt de arca et uniuersis bestiis terrae 9:11 statuam pactum meum 9:11 And I will establish my covenant with you, neither

shall all flesh be cut off any more by the waters of a flood; neither shall there any more be a flood to destroy the earth.

uobiscum et nequaquam ultra interficietur omnis caro aquis diluuii neque erit deinceps diluuium dissipans terram 9:12 9:12 And God said, This is the token of the covenant

which I make between me and you and every living creature that is with you, for perpetual generations:

dixitque Deus hoc signum foederis quod do inter me et uos et ad omnem animam uiuentem quae est uobiscum in generationes sempiternas 9:13 arcum meum ponam in nubibus et 9:13 I do set my bow in the cloud, and it shall be for a

token of a covenant between me and the earth.

erit signum foederis inter me et inter terram 9:14 cumque 9:14 And it shall come to pass, when I bring a cloud over

the earth, that the bow shall be seen in the cloud:

obduxero nubibus caelum apparebit arcus meus in nubibus 9:15 et recordabor foederis mei uobiscum et cum omni anima 9:15 And I will remember my covenant, which is between

me and you and every living creature of all flesh; and the waters shall no more become a flood to destroy all flesh.

uiuente quae carnem uegetat et non erunt ultra aquae diluuii ad delendam uniuersam carnem 9:16 eritque arcus in nubi- 9:16 And the bow shall be in the cloud; and I will look

upon it, that I may remember the everlasting covenant between God and every living creature of all flesh that is

bus et uidebo illum et recordabor foederis sempiterni quod upon the earth. pactum est inter Deum et inter omnem animam uiuentem uniuersae carnis quae est super terram 9:17 dixitque Deus 9:17 And God said unto Noah, This is the token of the

covenant, which I have established between me and all flesh that is upon the earth.

Noe hoc erit signum foederis quod constitui inter me et inter omnem carnem super terram 9:18 erant igitur filii Noe qui 9:18 And the sons of Noah, that went forth of the ark,

were Shem, and Ham, and Japheth: and Ham is the father of Canaan.

egressi sunt de arca Sem Ham et Iafeth porro Ham ipse est pater Chanaan 9:19 tres isti sunt filii Noe et ab his disse- 9:19 These are the three sons of Noah: and of them was

the whole earth overspread.

minatum est omne hominum genus super uniuersam terram 9:20 coepitque Noe uir agricola exercere terram et planta- 9:20 And Noah began to be an husbandman, and he

planted a vineyard:

uit uineam 9:21 bibensque uinum inebriatus est et nudatus 9:21 And he drank of the wine, and was drunken; and he

was uncovered within his tent.

9:22 And Ham, the father of Canaan, saw the nakedness in tabernaculo suo 9:22 quod cum uidisset Ham pater Cha-

of his father, and told his two brethren without.

naan uerenda scilicet patris sui esse nuda nuntiauit duobus

9:23 And Shem and Japheth took a garment, and laid it fratribus suis foras 9:23 at uero Sem et Iafeth pallium inpo-

upon both their shoulders, and went backward, and covered the nakedness of their father; and their faces were

backward, and they saw not their father’s nakedness. suerunt umeris suis et incedentes retrorsum operuerunt uerecunda patris sui faciesque eorum auersae erant et patris

9:24 And Noah awoke from his wine, and knew what his uirilia non uiderunt 9:24 euigilans autem Noe ex uino cum

younger son had done unto him.

9:25 And he said, Cursed be Canaan; a servant of servants didicisset quae fecerat ei filius suus minor 9:25 ait maledic-

shall he be unto his brethren.

9:26 And he said, Blessed be the LORD God of Shem; and tus Chanaan seruus seruorum erit fratribus suis 9:26 dixit-

Canaan shall be his servant.

que benedictus Dominus Deus Sem sit Chanaan seruus eius

9:27 God shall enlarge Japheth, and he shall dwell in the 9:27 dilatet Deus Iafeth et habitet in tabernaculis Sem sit-

tents of Shem; and Canaan shall be his servant. 9:28 And Noah lived after the flood three hundred and fifty que Chanaan seruus eius 9:28 uixit autem Noe post diluyears.

9:29 And all the days of Noah were nine hundred and fifty uium trecentis quinquaginta annis 9:29 et impleti sunt om-

years: and he died.

nes dies eius nongentorum quinquaginta annorum et mortuus est

10:1 Now these are the generations of the sons of Noah, 10:1 hae generationes filiorum Noe Sem Ham Iafeth nati-

Shem, Ham, and Japheth: and unto them were sons born after the flood.

que sunt eis filii post diluuium 10:2 filii Iafeth Gomer Ma-

10:2 The sons of Japheth; Gomer, and Magog, and Madai, and Javan, and Tubal, and Meshech, and Tiras. 10:3 And the sons of Gomer; Ashkenaz, and Riphath, and gog et Madai Iauan et Thubal et Mosoch et Thiras 10:3 Togarmah. 10:4 And the sons of Javan; Elishah, and Tarshish, Kittim, porro filii Gomer Aschenez et Rifath et Thogorma 10:4 filii and Dodanim. 10:5 By these were the isles of the Gentiles divided in their autem Iauan Elisa et Tharsis Cetthim et Dodanim 10:5 ab lands; every one after his tongue, after their families, in their nations.

his diuisae sunt insulae gentium in regionibus suis unusquis-

10:6 And the sons of Ham; Cush, and Mizraim, and Phut, que secundum linguam et familias in nationibus suis 10:6

and Canaan. 10:7 And the sons of Cush; Seba, and Havilah, and Sab- filii autem Ham Chus et Mesraim et Fut et Chanaan 10:7 tah, and Raamah, and Sabtechah: and the sons of Raamah; Sheba, and Dedan.

filii Chus Saba et Heuila et Sabatha et Regma et Sabathaca

10:8 And Cush begat Nimrod: he began to be a mighty filii Regma Saba et Dadan 10:8 porro Chus genuit Nemrod

one in the earth.

10:9 He was a mighty hunter before the LORD: wherefore ipse coepit esse potens in terra 10:9 et erat robustus uenator

it is said, Even as Nimrod the mighty hunter before the LORD.

coram Domino ab hoc exiuit prouerbium quasi Nemrod ro-

10:10 And the beginning of his kingdom was Babel, and bustus uenator coram Domino 10:10 fuit autem principium

Erech, and Accad, and Calneh, in the land of Shinar.

regni eius Babylon et Arach et Archad et Chalanne in terra

10:11 Out of that land went forth Asshur, and builded Nin- Sennaar 10:11 de terra illa egressus est Assur et aedificauit

eveh, and the city Rehoboth, and Calah,

10:12 And Resen between Nineveh and Calah: the same Nineuen et plateas ciuitatis et Chale 10:12 Resen quoque

is a great city.

10:13 And Mizraim begat Ludim, and Anamim, and inter Nineuen et Chale haec est ciuitas magna 10:13 at uero

Lehabim, and Naphtuhim,

Mesraim genuit Ludim et Anamim et Laabim Nepthuim

10:14 And Pathrusim, and Casluhim, (out of whom came 10:14 et Phetrusim et Cesluim de quibus egressi sunt Philis-

Philistim,) and Caphtorim.

Liber Genesis

thim et Capthurim 10:15 Chanaan autem genuit Sidonem 10:15 And Canaan begat Sidon his first born, and Heth,

primogenitum suum Ettheum 10:16 et Iebuseum et Amor- 10:16 And the Jebusite, and the Amorite, and the Girga-

site,

reum Gergeseum 10:17 Eueum et Araceum Sineum 10:18 10:17 And the Hivite, and the Arkite, and the Sinite, 10:18 And the Arvadite, and the Zemarite, and the Hamathite: and afterward were the families of the Canaanites spread abroad.

et Aradium Samariten et Amatheum et post haec disseminati sunt populi Chananeorum 10:19 factique sunt termini 10:19 And the border of the Canaanites was from Sidon,

as thou comest to Gerar, unto Gaza; as thou goest, unto Sodom, and Gomorrah, and Admah, and Zeboim, even

Chanaan uenientibus a Sidone Geraram usque Gazam donec unto Lasha. ingrediaris Sodomam et Gomorram et Adama et Seboim us-

que Lesa 10:20 hii filii Ham in cognationibus et linguis et 10:20 These are the sons of Ham, after their families, after

their tongues, in their countries, and in their nations.

generationibus terrisque et gentibus suis 10:21 de Sem quo- 10:21 Unto Shem also, the father of all the children of

Eber, the brother of Japheth the elder, even to him were children born.

que nati sunt patre omnium filiorum Eber fratre Iafeth maiore 10:22 filii Sem Aelam et Assur et Arfaxad et Lud et 10:22 The children of Shem; Elam, and Asshur, and

Arphaxad, and Lud, and Aram.

Aram 10:23 filii Aram Vs et Hul et Gether et Mes 10:24 at 10:23 And the children of Aram; Uz, and Hul, and Gether, and Mash. 10:24 And Arphaxad begat Salah; and Salah begat Eber.

uero Arfaxad genuit Sala de quo ortus est Eber 10:25 nati-

10:25 And unto Eber were born two sons: the name of one was Peleg; for in his days was the earth divided; and his

que sunt Eber filii duo nomen uni Faleg eo quod in diebus ^{brother’s name was Joktan.} eius diuisa sit terra et nomen fratris eius Iectan 10:26 qui 10:26 And Joktan begat Almodad, and Sheleph, and Haz-

armaveth, and Jerah,

Iectan genuit Helmodad et Saleph et Asarmoth Iare 10:27 10:27 And Hadoram, and Uzal, and Diklah, et Aduram et Vzal Decla 10:28 et Ebal et Abimahel Saba 10:28 And Obal, and Abimael, and Sheba, 10:29 et Ophir et Euila et Iobab omnes isti filii Iectan 10:30 10:29 And Ophir, and Havilah, and Jobab: all these were

the sons of Joktan.

10:30 And their dwelling was from Mesha, as thou goest unto Sephar a mount of the east.

et facta est habitatio eorum de Messa pergentibus usque Se-

phar montem orientalem 10:31 isti filii Sem secundum co- 10:31 These are the sons of Shem, after their families, af-

ter their tongues, in their lands, after their nations.

gnationes et linguas et regiones in gentibus suis 10:32 hae 10:32 These are the families of the sons of Noah, after

their generations, in their nations: and by these were the nations divided in the earth after the flood.

familiae Noe iuxta populos et nationes suas ab his diuisae sunt gentes in terra post diluuium

11:1 erat autem terra labii unius et sermonum eorundem 11:1 And the whole earth was of one language, and of one

speech.

11:2 cumque proficiscerentur de oriente inuenerunt campum 11:2 And it came to pass, as they journeyed from the east,

that they found a plain in the land of Shinar; and they dwelt there.

in terra Sennaar et habitauerunt in eo 11:3 dixitque alter

11:3 And they said one to another, Go to, let us make brick, and burn them thoroughly. And they had brick for stone, and slime had they for morter.

ad proximum suum uenite faciamus lateres et coquamus eos igni habueruntque lateres pro saxis et bitumen pro cemento 11:4 et dixerunt uenite faciamus nobis ciuitatem et turrem 11:4 And they said, Go to, let us build us a city and a

tower, whose top may reach unto heaven; and let us make us a name, lest we be scattered abroad upon the face of the

cuius culmen pertingat ad caelum et celebremus nomen no- whole earth. strum antequam diuidamur in uniuersas terras 11:5 descen- 11:5 And the LORD came down to see the city and the

tower, which the children of men builded.

dit autem Dominus ut uideret ciuitatem et turrem quam aedificabant filii Adam 11:6 et dixit ecce unus est populus et 11:6 And the LORD said, Behold, the people is one, and

they have all one language; and this they begin to do: and now nothing will be restrained from them, which they have imagined to do.

unum labium omnibus coeperuntque hoc facere nec desis-

11:7 Go to, let us go down, and there confound their lan- tent a cogitationibus suis donec eas opere conpleant 11:7

guage, that they may not understand one another’s speech.

uenite igitur descendamus et confundamus ibi linguam eo-

11:8 So the LORD scattered them abroad from thence rum ut non audiat unusquisque uocem proximi sui 11:8 at-

upon the face of all the earth: and they left off to build the city.

que ita diuisit eos Dominus ex illo loco in uniuersas terras

11:9 Therefore is the name of it called Babel; because the et cessauerunt aedificare ciuitatem 11:9 et idcirco uocatum

LORD did there confound the language of all the earth: and from thence did the LORD scatter them abroad upon

the face of all the earth. est nomen eius Babel quia ibi confusum est labium uniuersae terrae et inde dispersit eos Dominus super faciem cunc-

11:10 These are the generations of Shem: Shem was an tarum regionum 11:10 hae generationes Sem Sem centum

hundred years old, and begat Arphaxad two years after the flood:

erat annorum quando genuit Arfaxad biennio post diluuium

11:11 And Shem lived after he begat Arphaxad five hun- 11:11 uixitque Sem postquam genuit Arfaxad quingentos an-

dred years, and begat sons and daughters.

11:12 And Arphaxad lived five and thirty years, and begat nos et genuit filios et filias 11:12 porro Arfaxad uixit tri-

Salah:

11:13 And Arphaxad lived after he begat Salah four hun- ginta quinque annos et genuit Sale 11:13 uixitque Arfaxad

dred and three years, and begat sons and daughters.

postquam genuit Sale trecentis tribus annis et genuit filios et

11:14 And Salah lived thirty years, and begat Eber: filias 11:14 Sale quoque uixit triginta annis et genuit Eber 11:15 And Salah lived after he begat Eber four hundred 11:15 uixitque Sale postquam genuit Eber quadringentis tri-

and three years, and begat sons and daughters. 11:16 And Eber lived four and thirty years, and begat Pe- bus annis et genuit filios et filias 11:16 uixit autem Eber leg:

11:17 And Eber lived after he begat Peleg four hundred triginta quattuor annis et genuit Faleg 11:17 et uixit Eber

and thirty years, and begat sons and daughters.

postquam genuit Faleg quadringentis triginta annis et genuit

11:18 And Peleg lived thirty years, and begat Reu: filios et filias 11:18 uixit quoque Faleg triginta annis et ge-11:19 And Peleg lived after he begat Reu two hundred and nuit Reu 11:19 uixitque Faleg postquam genuit Reu ducen-

nine years, and begat sons and daughters.

11:20 And Reu lived two and thirty years, and begat tis nouem annis et genuit filios et filias 11:20 uixit autem

Serug:

11:21 And Reu lived after he begat Serug two hundred and Reu triginta duobus annis et genuit Sarug 11:21 uixitque

seven years, and begat sons and daughters.

Reu postquam genuit Sarug ducentis septem annis et genuit

11:22 And Serug lived thirty years, and begat Nahor: filios et filias 11:22 uixit uero Sarug triginta annis et genuit 11:23 And Serug lived after he begat Nahor two hundred Nahor 11:23 uixitque Sarug postquam genuit Nahor ducen-

years, and begat sons and daughters.

11:24 And Nahor lived nine and twenty years, and begat tos annos et genuit filios et filias 11:24 uixit autem Nahor

Terah:

11:25 And Nahor lived after he begat Terah an hundred uiginti nouem annis et genuit Thare 11:25 uixitque Nahor

and nineteen years, and begat sons and daughters.

postquam genuit Thare centum decem et nouem annos et ge-

11:26 And Terah lived seventy years, and begat Abram, nuit filios et filias 11:26 uixitque Thare septuaginta annis et

Nahor, and Haran. 11:27 Now these are the generations of Terah: Terah begat genuit Abram et Nahor et Aran 11:27 hae sunt autem gene-Abram, Nahor, and Haran; and Haran begat Lot.

rationes Thare Thare genuit Abram et Nahor et Aran porro

11:28 And Haran died before his father Terah in the land Aran genuit Loth 11:28 mortuusque est Aran ante Thare pa-

of his nativity, in Ur of the Chaldees.

Liber Genesis

trem suum in terra natiuitatis suae in Vr Chaldeorum 11:29 11:29 And Abram and Nahor took them wives: the name

of Abram’s wife was Sarai; and the name of Nahor’s wife, Milcah, the daughter of Haran, the father of Milcah, and

duxerunt autem Abram et Nahor uxores nomen autem uxoris the father of Iscah. Abram Sarai et nomen uxoris Nahor Melcha filia Aran patris Melchae et patris Ieschae 11:30 erat autem Sarai sterilis nec 11:30 But Sarai was barren; she had no child. habebat liberos 11:31 tulit itaque Thare Abram filium suum 11:31 And Terah took Abram his son, and Lot the son of

Haran his son’s son, and Sarai his daughter in law, his son Abram’s wife; and they went forth with them from Ur of the Chaldees, to go into the land of Canaan; and they came unto Haran, and dwelt there.

et Loth filium Aran filium filii sui et Sarai nurum suam uxorem Abram filii sui et eduxit eos de Vr Chaldeorum ut irent in terram Chanaan ueneruntque usque Haran et habitauerunt ibi 11:32 et facti sunt dies Thare ducentorum quinque an- 11:32 And the days of Terah were two hundred and five

years: and Terah died in Haran.

norum et mortuus est in Haran

12:1 dixit autem Dominus ad Abram egredere de terra tua 12:1 Now the LORD had said unto Abram, Get thee out

of thy country, and from thy kindred, and from thy father’s house, unto a land that I will shew thee:

et de cognatione tua et de domo patris tui in terram quam monstrabo tibi 12:2 faciamque te in gentem magnam et be- 12:2 And I will make of thee a great nation, and I will

bless thee, and make thy name great; and thou shalt be a blessing:

nedicam tibi et magnificabo nomen tuum erisque benedictus 12:3 benedicam benedicentibus tibi et maledicam maledicen- 12:3 And I will bless them that bless thee, and curse him

that curseth thee: and in thee shall all families of the earth be blessed.

tibus tibi atque in te benedicentur uniuersae cognationes terrae 12:4 egressus est itaque Abram sicut praeceperat ei Do- 12:4 So Abram departed, as the LORD had spoken unto

him; and Lot went with him: and Abram was seventy and five years old when he departed out of Haran.

minus et iuit cum eo Loth septuaginta quinque annorum erat Abram cum egrederetur de Haran 12:5 tulitque Sarai uxo- 12:5 And Abram took Sarai his wife, and Lot his brother’s

son, and all their substance that they had gathered, and the souls that they had gotten in Haran; and they went forth to go into the land of Canaan; and into the land of Canaan they came.

rem suam et Loth filium fratris sui uniuersamque substantiam quam possederant et animas quas fecerant in Haran et egressi sunt ut irent in terram Chanaan cumque uenissent in

eam 12:6 pertransiuit Abram terram usque ad locum Sy- 12:6 And Abram passed through the land unto the place of Sichem, unto the plain of Moreh. And the Canaanite was then in the land.

chem usque ad conuallem Inlustrem Chananeus autem tunc erat in terra 12:7 apparuitque Dominus Abram et dixit ei 12:7 And the LORD appeared unto Abram, and said, Unto

thy seed will I give this land: and there builded he an altar unto the LORD, who appeared unto him.

semini tuo dabo terram hanc qui aedificauit ibi altare Domino qui apparuerat ei 12:8 et inde transgrediens ad mon- 12:8 And he removed from thence unto a mountain on the

east of Bethel, and pitched his tent, having Bethel on the west, and Hai on the east: and there he builded an altar

tem qui erat contra orientem Bethel tetendit ibi tabernacu- unto the LORD, and called upon the name of the LORD. lum suum ab occidente habens Bethel et ab oriente Ai aedificauit quoque ibi altare Domino et inuocauit nomen eius 12:9 perrexitque Abram uadens et ultra progrediens ad me- 12:9 And Abram journeyed, going on still toward the

south.

ridiem 12:10 facta est autem fames in terra descenditque 12:10 And there was a famine in the land: and Abram

went down into Egypt to sojourn there; for the famine was grievous in the land.

Abram in Aegyptum ut peregrinaretur ibi praeualuerat enim

Biblia Sacra Vulgata

12:11 And it came to pass, when he was come near to enter fames in terra 12:11 cumque prope esset ut ingrederetur Ae-

into Egypt, that he said unto Sarai his wife, Behold now, I know that thou art a fair woman to look upon:

gyptum dixit Sarai uxori suae noui quod pulchra sis mulier

12:12 Therefore it shall come to pass, when the Egyptians 12:12 et quod cum uiderint te Aegyptii dicturi sunt uxor ip-

shall see thee, that they shall say, This is his wife: and they will kill me, but they will save thee alive.

sius est et interficient me et te reseruabunt 12:13 dic ergo

12:13 Say, I pray thee, thou art my sister: that it may be well with me for thy sake; and my soul shall live because of thee.

obsecro te quod soror mea sis ut bene sit mihi propter te et

12:14 And it came to pass, that, when Abram was come uiuat anima mea ob gratiam tui 12:14 cum itaque ingressus

into Egypt, the Egyptians beheld the woman that she was very fair.

esset Abram Aegyptum uiderunt Aegyptii mulierem quod

12:15 The princes also of Pharaoh saw her, and com- esset pulchra nimis 12:15 et nuntiauerunt principes Phara-

mended her before Pharaoh: and the woman was taken into Pharaoh’s house.

oni et laudauerunt eam apud illum et sublata est mulier in

12:16 And he entreated Abram well for her sake: and he domum Pharaonis 12:16 Abram uero bene usi sunt propter

had sheep, and oxen, and he asses, and menservants, and maidservants, and she asses, and camels.

illam fueruntque ei oues et boues et asini et serui et famulae

12:17 And the LORD plagued Pharaoh and his house with et asinae et cameli 12:17 flagellauit autem Dominus Phara-

great plagues because of Sarai Abram’s wife.

onem plagis maximis et domum eius propter Sarai uxorem

12:18 And Pharaoh called Abram and said, What is this Abram 12:18 uocauitque Pharao Abram et dixit ei quidnam

that thou hast done unto me? why didst thou not tell me that she was thy wife?

est quod fecisti mihi quare non indicasti quod uxor tua esset

12:19 Why saidst thou, She is my sister? so I might have 12:19 quam ob causam dixisti esse sororem tuam ut tollerem

taken her to me to wife: now therefore behold thy wife, take her, and go thy way.

eam mihi in uxorem nunc igitur ecce coniux tua accipe eam

12:20 And Pharaoh commanded his men concerning him: et uade 12:20 praecepitque Pharao super Abram uiris et de-

and they sent him away, and his wife, and all that he had.

duxerunt eum et uxorem illius et omnia quae habebat

13:1 And Abram went up out of Egypt, he, and his wife, 13:1 ascendit ergo Abram de Aegypto ipse et uxor eius et

and all that he had, and Lot with him, into the south.

omnia quae habebat et Loth cum eo ad australem plagam

13:2 And Abram was very rich in cattle, in silver, and in 13:2 erat autem diues ualde in possessione argenti et auri

gold.

13:3 And he went on his journeys from the south even 13:3 reuersusque est per iter quo uenerat a meridie in Bethel

to Bethel, unto the place where his tent had been at the beginning, between Bethel and Hai;

usque ad locum ubi prius fixerat tabernaculum inter Bethel

13:4 Unto the place of the altar, which he had make there et Ai 13:4 in loco altaris quod fecerat prius et inuocauit ibi

at the first: and there Abram called on the name of the LORD.

nomen Domini 13:5 sed et Loth qui erat cum Abram fue-

13:5 And Lot also, which went with Abram, had flocks, and herds, and tents.

13:6 And the land was not able to bear them, that they runt greges ouium et armenta et tabernacula 13:6 nec pote-

might dwell together: for their substance was great, so that they could not dwell together.

rat eos capere terra ut habitarent simul erat quippe substantia

13:7 And there was a strife between the herdmen of eorum multa et non quibant habitare communiter 13:7 unde

Abram’s cattle and the herdmen of Lot’s cattle: and the Canaanite and the Perizzite dwelled then in the land.

et facta est rixa inter pastores gregum Abram et Loth eo autem tempore Chananeus et Ferezeus habitabant in illa terra

13:8 And Abram said unto Lot, Let there be no strife, I 13:8 dixit ergo Abram ad Loth ne quaeso sit iurgium inter

pray thee, between me and thee, and between my herdmen and thy herdmen; for we be brethren.

me et te et inter pastores meos et pastores tuos fratres enim

Liber Genesis

sumus 13:9 ecce uniuersa terra coram te est recede a me ob- 13:9 Is not the whole land before thee? separate thyself, I pray thee, from me: if thou wilt take the left hand, then I will go to the right; or if thou depart to the right hand, then

secro si ad sinistram ieris ego ad dexteram tenebo si tu dexte- I will go to the left. ram elegeris ego ad sinistram pergam 13:10 eleuatis itaque 13:10 And Lot lifted up his eyes, and beheld all the plain

of Jordan, that it was well watered every where, before the LORD destroyed Sodom and Gomorrah, even as the garden of the LORD, like the land of Egypt, as thou comest unto Zoar.

Loth oculis uidit omnem circa regionem Iordanis quae uniuersa inrigabatur antequam subuerteret Dominus Sodomam et Gomorram sicut paradisus Domini et sicut Aegyptus uenientibus in Segor 13:11 elegitque sibi Loth regionem circa 13:11 Then Lot chose him all the plain of Jordan; and

Lot journeyed east: and they separated themselves the one from the other.

Iordanem et recessit ab oriente diuisique sunt alterutrum a fratre suo 13:12 Abram habitauit in terra Chanaan Loth mo- 13:12 Abram dwelled in the land of Canaan, and Lot

dwelled in the cities of the plain, and pitched his tent toward Sodom.

ratus est in oppidis quae erant circa Iordanem et habitauit in Sodomis 13:13 homines autem Sodomitae pessimi erant et 13:13 But the men of Sodom were wicked and sinners be-

fore the LORD exceedingly.

peccatores coram Domino nimis 13:14 dixitque Dominus 13:14 And the LORD said unto Abram, after that Lot was

separated from him, Lift up now thine eyes, and look from the place where thou art northward, and southward, and

ad Abram postquam diuisus est Loth ab eo leua oculos tuos eastward, and westward: et uide a loco in quo nunc es ad aquilonem et ad meridiem ad

orientem et ad occidentem 13:15 omnem terram quam con- 13:15 For all the land which thou seest, to thee will I give

it, and to thy seed for ever.

spicis tibi dabo et semini tuo usque in sempiternum 13:16 13:16 And I will make thy seed as the dust of the earth: so

that if a man can number the dust of the earth, then shall thy seed also be numbered.

faciamque semen tuum sicut puluerem terrae si quis potest hominum numerare puluerem semen quoque tuum numerare poterit 13:17 surge et perambula terram in longitudine et in 13:17 Arise, walk through the land in the length of it and

in the breadth of it; for I will give it unto thee.

latitudine sua quia tibi daturus sum eam 13:18 mouens igi- 13:18 Then Abram removed his tent, and came and dwelt

in the plain of Mamre, which is in Hebron, and built there an altar unto the LORD.

tur Abram tabernaculum suum uenit et habitauit iuxta conuallem Mambre quod est in Hebron aedificauitque ibi altare Domino

14:1 factum est autem in illo tempore ut Amrafel rex Sen- 14:1 And it came to pass in the days of Amraphel king

of Shinar, Arioch king of Ellasar, Chedorlaomer king of Elam, and Tidal king of nations;

naar et Arioch rex Ponti et Chodorlahomor rex Aelamitarum et Thadal rex Gentium 14:2 inirent bellum contra Bara re- 14:2 That these made war with Bera king of Sodom, and

with Birsha king of Gomorrah, Shinab king of Admah, and Shemeber king of Zeboiim, and the king of Bela, which is

gem Sodomorum et contra Bersa regem Gomorrae et con- Zoar. tra Sennaab regem Adamae et contra Semeber regem Se-

boim contraque regem Balae ipsa est Segor 14:3 omnes hii 14:3 All these were joined together in the vale of Siddim,

which is the salt sea.

conuenerunt in uallem Siluestrem quae nunc est mare Salis 14:4 duodecim enim annis seruierant Chodorlahomor et ter- 14:4 Twelve years they served Chedorlaomer, and in the

thirteenth year they rebelled.

tiodecimo anno recesserunt ab eo 14:5 igitur anno quarto- 14:5 And in the fourteenth year came Chedorlaomer, and

the kings that were with him, and smote the Rephaims in Ashteroth Karnaim, and the Zuzims in Ham, and the

decimo uenit Chodorlahomor et reges qui erant cum eo per- Emins in Shaveh Kiriathaim,

Biblia Sacra Vulgata

cusseruntque Rafaim in Astharothcarnaim et Zuzim cum eis

14:6 And the Horites in their mount Seir, unto Elparan, et Emim in Sauecariathaim 14:6 et Chorreos in montibus which is by the wilderness.

14:7 And they returned, and came to Enmishpat, which is Seir usque ad campestria Pharan quae est in solitudine 14:7

Kadesh, and smote all the country of the Amalekites, and also the Amorites, that dwelt in Hazezontamar.

reuersique sunt et uenerunt ad fontem Mesfat ipsa est Cades et percusserunt omnem regionem Amalechitarum et Amor-

14:8 And there went out the king of Sodom, and the king reum qui habitabat in Asasonthamar 14:8 et egressi sunt rex

of Gomorrah, and the king of Admah, and the king of Zeboiim, and the king of Bela (the same is Zoar;) and they

joined battle with them in the vale of Siddim; Sodomorum et rex Gomorrae rexque Adamae et rex Seboim

necnon et rex Balae quae est Segor et direxerunt contra eos

14:9 With Chedorlaomer the king of Elam, and with Tidal aciem in ualle Siluestri 14:9 scilicet aduersum Chodorlaho-

king of nations, and Amraphel king of Shinar, and Arioch king of Ellasar; four kings with five.

mor regem Aelamitarum et Thadal regem Gentium et Amrafel regem Sennaar et Arioch regem Ponti quattuor reges

14:10 And the vale of Siddim was full of slimepits; and aduersus quinque 14:10 uallis autem Siluestris habebat pu-

the kings of Sodom and Gomorrah fled, and fell there; and they that remained fled to the mountain.

teos multos bituminis itaque rex Sodomorum et Gomorrae terga uerterunt cecideruntque ibi et qui remanserant fuge-

14:11 And they took all the goods of Sodom and Gomor- runt ad montem 14:11 tulerunt autem omnem substantiam

rah, and all their victuals, and went their way.

Sodomorum et Gomorrae et uniuersa quae ad cibum perti-

14:12 And they took Lot, Abram’s brother’s son, who nent et abierunt 14:12 necnon et Loth et substantiam eius

dwelt in Sodom, and his goods, and departed.

14:13 And there came one that had escaped, and told filium fratris Abram qui habitabat in Sodomis 14:13 et ecce

Abram the Hebrew; for he dwelt in the plain of Mamre the Amorite, brother of Eshcol, and brother of Aner: and

these were confederate with Abram. unus qui euaserat nuntiauit Abram Hebraeo qui habitabat in

conualle Mambre Amorrei fratris Eschol et fratris Aner hii

14:14 And when Abram heard that his brother was taken enim pepigerant foedus cum Abram 14:14 quod cum audis-

captive, he armed his trained servants, born in his own house, three hundred and eighteen, and pursued them unto

Dan. set Abram captum uidelicet Loth fratrem suum numerauit

expeditos uernaculos suos trecentos decem et octo et per-

14:15 And he divided himself against them, he and his secutus est eos usque Dan 14:15 et diuisis sociis inruit su-

servants, by night, and smote them, and pursued them unto Hobah, which is on the left hand of Damascus.

per eos nocte percussitque eos et persecutus est usque Hoba

14:16 And he brought back all the goods, and also brought quae est ad leuam Damasci 14:16 reduxitque omnem subsagain his brother Lot, and his goods, and the women also, and the people.

tantiam et Loth fratrem suum cum substantia illius mulieres

14:17 And the king of Sodom went out to meet him after quoque et populum 14:17 egressus est autem rex Sodomo-

his return from the slaughter of Chedorlaomer, and of the kings that were with him, at the valley of Shaveh, which is

the king’s dale. rum in occursum eius postquam reuersus est a caede Cho-

dorlahomor et regum qui cum eo erant in ualle Saue quae est

14:18 And Melchizedek king of Salem brought forth bread uallis Regis 14:18 at uero Melchisedech rex Salem profe-

and wine: and he was the priest of the most high God.

14:19 And he blessed him, and said, Blessed be Abram of rens panem et uinum erat enim sacerdos Dei altissimi 14:19

the most high God, possessor of heaven and earth:

benedixit ei et ait benedictus Abram Deo excelso qui crea-

Liber Genesis

uit caelum et terram 14:20 et benedictus Deus excelsus quo 14:20 And blessed be the most high God, which hath de-

livered thine enemies into thy hand. And he gave him tithes of all.

protegente hostes in manibus tuis sunt et dedit ei decimas ex omnibus 14:21 dixit autem rex Sodomorum ad Abram 14:21 And the king of Sodom said unto Abram, Give me

the persons, and take the goods to thyself.

da mihi animas cetera tolle tibi 14:22 qui respondit ei leuo 14:22 And Abram said to the king of Sodom, I have lift up

mine hand unto the LORD, the most high God, the possessor of heaven and earth,

manum meam ad Dominum Deum excelsum possessorem caeli et terrae 14:23 quod a filo subteminis usque ad cor- 14:23 That I will not take from a thread even to a shoe-

latchet, and that I will not take any thing that is thine, lest thou shouldest say, I have made Abram rich:

rigiam caligae non accipiam ex omnibus quae tua sunt ne dicas ego ditaui Abram 14:24 exceptis his quae comede- 14:24 Save only that which the young men have eaten, and

the portion of the men which went with me, Aner, Eshcol, and Mamre; let them take their portion.

runt iuuenes et partibus uirorum qui uenerunt mecum Aner Eschol et Mambre isti accipient partes suas

15:1 his itaque transactis factus est sermo Domini ad Ab- 15:1 After these things the word of the LORD came unto

Abram in a vision, saying, Fear not, Abram: I am thy shield, and thy exceeding great reward.

ram per uisionem dicens noli timere Abram ego protector tuus sum et merces tua magna nimis 15:2 dixitque Abram 15:2 And Abram said, LORD God, what wilt thou give

me, seeing I go childless, and the steward of my house is this Eliezer of Damascus?

Domine Deus quid dabis mihi ego uadam absque liberis et filius procuratoris domus meae iste Damascus Eliezer 15:3 15:3 And Abram said, Behold, to me thou hast given no

seed: and, lo, one born in my house is mine heir.

addiditque Abram mihi autem non dedisti semen et ecce uernaculus meus heres meus erit 15:4 statimque sermo Domini 15:4 And, behold, the word of the LORD came unto him,

saying, This shall not be thine heir; but he that shall come forth out of thine own bowels shall be thine heir.

factus est ad eum dicens non erit hic heres tuus sed qui egredietur de utero tuo ipsum habebis heredem 15:5 eduxit- 15:5 And he brought him forth abroad, and said, Look now

toward heaven, and tell the stars, if thou be able to number them: and he said unto him, So shall thy seed be.

que eum foras et ait illi suspice caelum et numera stellas si potes et dixit ei sic erit semen tuum 15:6 credidit Domino 15:6 And he believed in the LORD; and he counted it to

him for righteousness.

et reputatum est ei ad iustitiam 15:7 dixitque ad eum ego 15:7 And he said unto him, I am the LORD that brought

thee out of Ur of the Chaldees, to give thee this land to inherit it.

Dominus qui eduxi te de Vr Chaldeorum ut darem tibi terram istam et possideres eam 15:8 at ille ait Domine Deus 15:8 And he said, LORD God, whereby shall I know that

I shall inherit it?

unde scire possum quod possessurus sim eam 15:9 respon- 15:9 And he said unto him, Take me an heifer of three

years old, and a she goat of three years old, and a ram of three years old, and a turtledove, and a young pigeon.

dens Dominus sume inquit mihi uaccam triennem et capram trimam et arietem annorum trium turturem quoque et columbam 15:10 qui tollens uniuersa haec diuisit per me- 15:10 And he took unto him all these, and divided them in

the midst, and laid each piece one against another: but the birds divided he not.

dium et utrasque partes contra se altrinsecus posuit aues autem non diuisit 15:11 descenderuntque uolucres super ca- 15:11 And when the fowls came down upon the carcases,

Abram drove them away.

dauera et abigebat eas Abram 15:12 cumque sol occumbe- 15:12 And when the sun was going down, a deep sleep fell

upon Abram; and, lo, an horror of great darkness fell upon him.

ret sopor inruit super Abram et horror magnus et tenebrosus inuasit eum 15:13 dictumque est ad eum scito praenos- 15:13 And he said unto Abram, Know of a surety that thy

seed shall be a stranger in a land that is not their’s, and shall serve them; and they shall afflict them four hundred years;

Biblia Sacra Vulgata

cens quod peregrinum futurum sit semen tuum in terra non sua et subicient eos seruituti et adfligent quadringentis annis

15:14 And also that nation, whom they shall serve, will I 15:14 uerumtamen gentem cui seruituri sunt ego iudicabo et

judge: and afterward shall they come out with great substance.

post haec egredientur cum magna substantia 15:15 tu autem

15:15 And thou shalt go to thy fathers in peace; thou shalt be buried in a good old age.

15:16 But in the fourth generation they shall come hither ibis ad patres tuos in pace sepultus in senectute bona 15:16

again: for the iniquity of the Amorites is not yet full.

generatione autem quarta reuertentur huc necdum enim conpletae sunt iniquitates Amorreorum usque ad praesens tem-

15:17 And it came to pass, that, when the sun went down, pus 15:17 cum ergo occubuisset sol facta est caligo tene-

and it was dark, behold a smoking furnace, and a burning lamp that passed between those pieces.

brosa et apparuit clibanus fumans et lampas ignis transiens

15:18 In the same day the LORD made a covenant with inter diuisiones illas 15:18 in die illo pepigit Dominus cum

Abram, saying, Unto thy seed have I given this land, from the river of Egypt unto the great river, the river Euphrates:

Abram foedus dicens semini tuo dabo terram hanc a fluuio

15:19 The Kenites, and the Kenizzites, and the Kad- Aegypti usque ad fluuium magnum flumen Eufraten 15:19

monites,

15:20 And the Hittites, and the Perizzites, and the Cineos et Cenezeos et Cedmoneos 15:20 et Hettheos et Fe-

Rephaims, 15:21 And the Amorites, and the Canaanites, and the Gir- rezeos Rafaim quoque 15:21 et Amorreos et Chananeos et gashites, and the Jebusites.

Gergeseos et Iebuseos

16:1 Now Sarai Abram’s wife bare him no children: and 16:1 igitur Sarai uxor Abram non genuerat liberos sed ha-

she had an handmaid, an Egyptian, whose name was Hagar.

bens ancillam aegyptiam nomine Agar 16:2 dixit marito

16:2 And Sarai said unto Abram, Behold now, the LORD hath restrained me from bearing: I pray thee, go in unto my maid; it may be that I may obtain children by her. And Abram hearkened to the voice of Sarai.

suo ecce conclusit me Dominus ne parerem ingredere ad ancillam meam si forte saltem ex illa suscipiam filios cum-

16:3 And Sarai Abram’s wife took Hagar her maid the que ille adquiesceret deprecanti 16:3 tulit Agar Aegyptiam

Egyptian, after Abram had dwelt ten years in the land of Canaan, and gave her to her husband Abram to be his wife.

ancillam suam post annos decem quam habitare coeperant

16:4 And he went in unto Hagar, and she conceived: and in terra Chanaan et dedit eam uiro suo uxorem 16:4 qui when she saw that she had conceived, her mistress was despised in her eyes.

ingressus est ad eam at illa concepisse se uidens despexit

16:5 And Sarai said unto Abram, My wrong be upon thee: dominam suam 16:5 dixitque Sarai ad Abram inique agis

I have given my maid into thy bosom; and when she saw that she had conceived, I was despised in her eyes: the

LORD judge between me and thee. contra me ego dedi ancillam meam in sinum tuum quae ui-

dens quod conceperit despectui me habet iudicet Dominus

16:6 But Abram said unto Sarai, Behold, thy maid is in inter me et te 16:6 cui respondens Abram ecce ait ancilla

thine hand; do to her as it pleaseth thee. And when Sarai dealt hardly with her, she fled from her face.

tua in manu tua est utere ea ut libet adfligente igitur eam

16:7 And the angel of the LORD found her by a fountain Sarai fugam iniit 16:7 cumque inuenisset illam angelus Do-

of water in the wilderness, by the fountain in the way to Shur.

mini iuxta fontem aquae in solitudine qui est in uia Sur 16:8

16:8 And he said, Hagar, Sarai’s maid, whence camest thou? and whither wilt thou go? And she said, I flee from the face of my mistress Sarai.

dixit ad eam Agar ancilla Sarai unde uenis et quo uadis quae

16:9 And the angel of the LORD said unto her, Return to respondit a facie Sarai dominae meae ego fugio 16:9 dixit-

thy mistress, and submit thyself under her hands.

que ei angelus Domini reuertere ad dominam tuam et humi-

Liber Genesis

liare sub manibus ipsius 16:10 et rursum multiplicans in- 16:10 And the angel of the LORD said unto her, I will

multiply thy seed exceedingly, that it shall not be numbered for multitude.

quit multiplicabo semen tuum et non numerabitur prae multitudine 16:11 ac deinceps ecce ait concepisti et paries fi- 16:11 And the angel of the LORD said unto her, Behold,

thou art with child and shalt bear a son, and shalt call his name Ishmael; because the LORD hath heard thy afflic-

lium uocabisque nomen eius Ismahel eo quod audierit Do- tion. minus adflictionem tuam 16:12 hic erit ferus homo manus 16:12 And he will be a wild man; his hand will be against

every man, and every man’s hand against him; and he shall dwell in the presence of all his brethren.

eius contra omnes et manus omnium contra eum et e regione uniuersorum fratrum suorum figet tabernacula 16:13 uoca- 16:13 And she called the name of the LORD that spake

unto her, Thou God seest me: for she said, Have I also here looked after him that seeth me?

uit autem nomen Domini qui loquebatur ad eam Tu Deus qui uidisti me dixit enim profecto hic uidi posteriora uidentis me 16:14 propterea appellauit puteum illum puteum Viuentis et 16:14 Wherefore the well was called Beerlahairoi; behold,

it is between Kadesh and Bered.

uidentis me ipse est inter Cades et Barad 16:15 peperitque 16:15 And Hagar bare Abram a son: and Abram called his

son’s name, which Hagar bare, Ishmael.

Abrae filium qui uocauit nomen eius Ismahel 16:16 octo- 16:16 And Abram was fourscore and six years old, when

Hagar bare Ishmael to Abram.

ginta et sex annorum erat quando peperit ei Agar Ismahelem

17:1 postquam uero nonaginta et nouem annorum esse 17:1 And when Abram was ninety years old and nine, the LORD appeared to Abram, and said unto him, I am the Almighty God; walk before me, and be thou perfect.

coeperat apparuit ei Dominus dixitque ad eum ego Deus omnipotens ambula coram me et esto perfectus 17:2 po- 17:2 And I will make my covenant between me and thee,

and will multiply thee exceedingly.

namque foedus meum inter me et te et multiplicabo te uehementer nimis 17:3 cecidit Abram pronus in faciem 17:4 17:3 And Abram fell on his face: and God talked with

him, saying,

17:4 As for me, behold, my covenant is with thee, and thou shalt be a father of many nations.

dixitque ei Deus ego sum et pactum meum tecum erisque

pater multarum gentium 17:5 nec ultra uocabitur nomen 17:5 Neither shall thy name any more be called Abram, but thy name shall be Abraham; for a father of many nations have I made thee.

tuum Abram sed appellaberis Abraham quia patrem multarum gentium constitui te 17:6 faciamque te crescere uehe- 17:6 And I will make thee exceeding fruitful, and I will

make nations of thee, and kings shall come out of thee.

mentissime et ponam in gentibus regesque ex te egredientur 17:7 et statuam pactum meum inter me et te et inter semen 17:7 And I will establish my covenant between me and

thee and thy seed after thee in their generations for an everlasting covenant, to be a God unto thee, and to thy seed

tuum post te in generationibus suis foedere sempiterno ut after thee. sim Deus tuus et seminis tui post te 17:8 daboque tibi et se- 17:8 And I will give unto thee, and to thy seed after thee,

the land wherein thou art a stranger, all the land of Canaan, for an everlasting possession; and I will be their God.

mini tuo terram peregrinationis tuae omnem terram Chanaan in possessionem aeternam eroque Deus eorum 17:9 dixit 17:9 And God said unto Abraham, Thou shalt keep my

covenant therefore, thou, and thy seed after thee in their generations.

iterum Deus ad Abraham et tu ergo custodies pactum meum et semen tuum post te in generationibus suis 17:10 hoc est 17:10 This is my covenant, which ye shall keep, between

me and you and thy seed after thee; Every man child among you shall be circumcised.

pactum quod obseruabitis inter me et uos et semen tuum post te circumcidetur ex uobis omne masculinum 17:11 et 17:11 And ye shall circumcise the flesh of your foreskin;

and it shall be a token of the covenant betwixt me and you.

circumcidetis carnem praeputii uestri ut sit in signum foe-

Biblia Sacra Vulgata

17:12 And he that is eight days old shall be circumcised deris inter me et uos 17:12 infans octo dierum circumcide-

among you, every man child in your generations, he that is born in the house, or bought with money of any stranger,

which is not of thy seed. tur in uobis omne masculinum in generationibus uestris tam

uernaculus quam empticius circumcidetur et quicumque non

17:13 He that is born in thy house, and he that is bought fuerit de stirpe uestra 17:13 eritque pactum meum in carne

with thy money, must needs be circumcised: and my covenant shall be in your flesh for an everlasting covenant.

uestra in foedus aeternum 17:14 masculus cuius praeputii

17:14 And the uncircumcised man child whose flesh of his foreskin is not circumcised, that soul shall be cut off from his people; he hath broken my covenant.

caro circumcisa non fuerit delebitur anima illa de populo suo

17:15 And God said unto Abraham, As for Sarai thy wife, quia pactum meum irritum fecit 17:15 dixit quoque Deus ad

thou shalt not call her name Sarai, but Sarah shall her name be.

Abraham Sarai uxorem tuam non uocabis Sarai sed Sarram

17:16 And I will bless her, and give thee a son also of her: 17:16 et benedicam ei et ex illa dabo tibi filium cui benedic-

yea, I will bless her, and she shall be a mother of nations; kings of people shall be of her.

turus sum eritque in nationes et reges populorum orientur ex

17:17 Then Abraham fell upon his face, and laughed, and eo 17:17 cecidit Abraham in faciem et risit dicens in corde

said in his heart, Shall a child be born unto him that is an hundred years old? and shall Sarah, that is ninety years

old, bear? suo putasne centenario nascetur filius et Sarra nonagenaria 17:18 And Abraham said unto God, O that Ishmael might pariet 17:18 dixitque ad Deum utinam Ismahel uiuat coram

live before thee!

17:19 And God said, Sarah thy wife shall bear thee a son te 17:19 et ait Deus ad Abraham Sarra uxor tua pariet tibi

indeed; and thou shalt call his name Isaac: and I will establish my covenant with him for an everlasting covenant,

and with his seed after him. filium uocabisque nomen eius Isaac et constituam pactum

meum illi in foedus sempiternum et semini eius post eum

17:20 And as for Ishmael, I have heard thee: Behold, I 17:20 super Ismahel quoque exaudiui te ecce benedicam ei et

have blessed him, and will make him fruitful, and will multiply him exceedingly; twelve princes shall he beget,

and I will make him a great nation. augebo et multiplicabo eum ualde duodecim duces genera-

17:21 But my covenant will I establish with Isaac, which bit et faciam illum in gentem magnam 17:21 pactum uero Sarah shall bear unto thee at this set time in the next year.

meum statuam ad Isaac quem pariet tibi Sarra tempore isto

17:22 And he left off talking with him, and God went up in anno altero 17:22 cumque finitus esset sermo loquentis

from Abraham.

17:23 And Abraham took Ishmael his son, and all that cum eo ascendit Deus ab Abraham 17:23 tulit autem Ab-

were born in his house, and all that were bought with his money, every male among the men of Abraham’s house; and circumcised the flesh of their foreskin in the selfsame day, as God had said unto him.

raham Ismahelem filium suum et omnes uernaculos domus suae uniuersosque quos emerat cunctos mares ex omnibus uiris domus suae et circumcidit carnem praeputii eorum sta-

17:24 And Abraham was ninety years old and nine, when tim in ipsa die sicut praeceperat ei Deus 17:24 nonaginta

he was circumcised in the flesh of his foreskin.

nouem erat annorum quando circumcidit carnem praeputii

17:25 And Ishmael his son was thirteen years old, when sui 17:25 et Ismahel filius eius tredecim annos impleuerat

he was circumcised in the flesh of his foreskin.

17:26 In the selfsame day was Abraham circumcised, and tempore circumcisionis suae 17:26 eadem die circumcisus

Ishmael his son.

17:27 And all the men of his house, born in the house, and est Abraham et Ismahel filius eius 17:27 et omnes uiri do-

bought with money of the stranger, were circumcised with him.

mus illius tam uernaculi quam empticii et alienigenae pariter circumcisi sunt

18:1 And the LORD appeared unto him in the plains of 18:1 apparuit autem ei Dominus in conualle Mambre se-

Mamre: and he sat in the tent door in the heat of the day;

Liber Genesis

denti in ostio tabernaculi sui in ipso feruore diei 18:2 cum- 18:2 And he lift up his eyes and looked, and, lo, three men

stood by him: and when he saw them, he ran to meet them from the tent door, and bowed himself toward the ground,

que eleuasset oculos apparuerunt ei tres uiri stantes propter eum quos cum uidisset cucurrit in occursum eorum de ostio tabernaculi et adorauit in terra 18:3 et dixit Domine si 18:3 And said, My LORD, if now I have found favour in

thy sight, pass not away, I pray thee, from thy servant:

inueni gratiam in oculis tuis ne transeas seruum tuum 18:4 18:4 Let a little water, I pray you, be fetched, and wash

your feet, and rest yourselves under the tree:

sed adferam pauxillum aquae et lauate pedes uestros et requiescite sub arbore 18:5 ponam buccellam panis et confor- 18:5 And I will fetch a morsel of bread, and comfort ye

your hearts; after that ye shall pass on: for therefore are ye come to your servant. And they said, So do, as thou hast

tate cor uestrum postea transibitis idcirco enim declinastis ad said. seruum uestrum qui dixerunt fac ut locutus es 18:6 festina- 18:6 And Abraham hastened into the tent unto Sarah, and

said, Make ready quickly three measures of fine meal, knead it, and make cakes upon the hearth.

uit Abraham in tabernaculum ad Sarram dixitque ei adcelera tria sata similae commisce et fac subcinericios panes 18:7 18:7 And Abraham ran unto the herd, and fetcht a calf

tender and good, and gave it unto a young man; and he hasted to dress it.

ipse uero ad armentum cucurrit et tulit inde uitulum tenerrimum et optimum deditque puero qui festinauit et coxit illum 18:8 tulit quoque butyrum et lac et uitulum quem coxerat et 18:8 And he took butter, and milk, and the calf which he

had dressed, and set it before them; and he stood by them under the tree, and they did eat.

posuit coram eis ipse uero stabat iuxta eos sub arbore 18:9

18:9 And they said unto him, Where is Sarah thy wife? And he said, Behold, in the tent.

cumque comedissent dixerunt ad eum ubi est Sarra uxor tua ille respondit ecce in tabernaculo est 18:10 cui dixit reuer- 18:10 And he said, I will certainly return unto thee accord-

ing to the time of life; and, lo, Sarah thy wife shall have a son. And Sarah heard it in the tent door, which was behind

tens ueniam ad te tempore isto uita comite et habebit filium him. Sarra uxor tua quo audito Sarra risit post ostium tabernaculi 18:11 erant autem ambo senes prouectaeque aetatis et 18:11 Now Abraham and Sarah were old and well stricken

in age; and it ceased to be with Sarah after the manner of women.

desierant Sarrae fieri muliebria 18:12 quae risit occulte di-

18:12 Therefore Sarah laughed within herself, saying, After I am waxed old shall I have pleasure, my lord being old also?

cens postquam consenui et dominus meus uetulus est uoluptati operam dabo 18:13 dixit autem Dominus ad Abraham 18:13 And the LORD said unto Abraham, Wherefore did

Sarah laugh, saying, Shall I of a surety bear a child, which am old?

quare risit Sarra dicens num uere paritura sum anus 18:14

18:14 Is any thing too hard for the LORD? At the time appointed I will return unto thee, according to the time of life, and Sarah shall have a son.

numquid Deo est quicquam difficile iuxta condictum reuertar ad te hoc eodem tempore uita comite et habebit Sarra filium 18:15 negauit Sarra dicens non risi timore perter- 18:15 Then Sarah denied, saying, I laughed not; for she

was afraid. And he said, Nay; but thou didst laugh.

rita Dominus autem non est inquit ita sed risisti 18:16 cum 18:16 And the men rose up from thence, and looked to-

ward Sodom: and Abraham went with them to bring them on the way.

ergo surrexissent inde uiri direxerunt oculos suos contra Sodomam et Abraham simul gradiebatur deducens eos 18:17 18:17 And the LORD said, Shall I hide from Abraham that

thing which I do;

dixitque Dominus num celare potero Abraham quae gesturus sum 18:18 cum futurus sit in gentem magnam ac robu- 18:18 Seeing that Abraham shall surely become a great

and mighty nation, and all the nations of the earth shall be blessed in him?

stissimam et benedicendae sint in illo omnes nationes ter-

Biblia Sacra Vulgata

18:19 For I know him, that he will command his children rae 18:19 scio enim quod praecepturus sit filiis suis et do-

and his household after him, and they shall keep the way of the LORD, to do justice and judgment; that the LORD may bring upon Abraham that which he hath spoken of him.

mui suae post se ut custodiant uiam Domini et faciant iustitiam et iudicium ut adducat Dominus propter Abraham

18:20 And the LORD said, Because the cry of Sodom and omnia quae locutus est ad eum 18:20 dixit itaque Dominus

Gomorrah is great, and because their sin is very grievous;

clamor Sodomorum et Gomorrae multiplicatus est et pec-

18:21 I will go down now, and see whether they have done catum earum adgrauatum est nimis 18:21 descendam et ui-

altogether according to the cry of it, which is come unto me; and if not, I will know.

debo utrum clamorem qui uenit ad me opere conpleuerint an

18:22 And the men turned their faces from thence, and non est ita ut sciam 18:22 conuerteruntque se inde et abie-

went toward Sodom: but Abraham stood yet before the LORD.

runt Sodomam Abraham uero adhuc stabat coram Domino

18:23 And Abraham drew near, and said, Wilt thou also 18:23 et adpropinquans ait numquid perdes iustum cum im-

destroy the righteous with the wicked?

18:24 Peradventure there be fifty righteous within the city: pio 18:24 si fuerint quinquaginta iusti in ciuitate peribunt

wilt thou also destroy and not spare the place for the fifty righteous that are therein?

simul et non parces loco illi propter quinquaginta iustos si

18:25 That be far from thee to do after this manner, to fuerint in eo 18:25 absit a te ut rem hanc facias et occi-

slay the righteous with the wicked: and that the righteous should be as the wicked, that be far from thee: Shall not

the Judge of all the earth do right? das iustum cum impio fiatque iustus sicut impius non est

hoc tuum qui iudicas omnem terram nequaquam facies iu-

18:26 And the LORD said, If I find in Sodom fifty righ- dicium 18:26 dixitque Dominus ad eum si inuenero Sodo-

teous within the city, then I will spare all the place for their sakes.

mis quinquaginta iustos in medio ciuitatis dimittam omni

18:27 And Abraham answered and said, Behold now, I loco propter eos 18:27 respondens Abraham ait quia semel

have taken upon me to speak unto the LORD, which am but dust and ashes:

coepi loquar ad Dominum meum cum sim puluis et cinis

18:28 Peradventure there shall lack five of the fifty righ- 18:28 quid si minus quinquaginta iustis quinque fuerint de-

teous: wilt thou destroy all the city for lack of five? And he said, If I find there forty and five, I will not destroy it.

lebis propter quinque uniuersam urbem et ait non delebo si

18:29 And he spake unto him yet again, and said, Perad- inuenero ibi quadraginta quinque 18:29 rursumque locutus

venture there shall be forty found there. And he said, I will not do it for forty’s sake.

est ad eum sin autem quadraginta inuenti fuerint quid facies

18:30 And he said unto him, Oh let not the LORD be an- ait non percutiam propter quadraginta 18:30 ne quaeso in-

gry, and I will speak: Peradventure there shall thirty be found there. And he said, I will not do it, if I find thirty

there. quit indigneris Domine si loquar quid si inuenti fuerint ibi 18:31 And he said, Behold now, I have taken upon me to triginta respondit non faciam si inuenero ibi triginta 18:31

speak unto the LORD: Peradventure there shall be twenty found there. And he said, I will not destroy it for twenty’s

sake. quia semel ait coepi loquar ad Dominum meum quid si in-

uenti fuerint ibi uiginti dixit non interficiam propter uiginti

18:32 And he said, Oh let not the LORD be angry, and 18:32 obsecro inquit ne irascaris Domine si loquar adhuc se-

I will speak yet but this once: Peradventure ten shall be found there. And he said, I will not destroy it for ten’s

sake. mel quid si inuenti fuerint ibi decem dixit non delebo prop-18:33 And the LORD went his way, as soon as he had left ter decem 18:33 abiit Dominus postquam cessauit loqui ad

communing with Abraham: and Abraham returned unto his place.

Abraham et ille reuersus est in locum suum

19:1 And there came two angels to Sodom at even; and 19:1 ueneruntque duo angeli Sodomam uespere sedente

Lot sat in the gate of Sodom: and Lot seeing them rose up to meet them; and he bowed himself with his face toward the ground;

Liber Genesis

Loth in foribus ciuitatis qui cum uidisset surrexit et iuit obuiam eis adorauitque pronus in terra 19:2 et dixit obsecro 19:2 And he said, Behold now, my lords, turn in, I pray

you, into your servant’s house, and tarry all night, and wash your feet, and ye shall rise up early, and go on your ways. And they said, Nay; but we will abide in the street all night.

domini declinate in domum pueri uestri et manete ibi lauate pedes uestros et mane proficiscimini in uiam uestram qui dixerunt minime sed in platea manebimus 19:3 conpulit il- 19:3 And he pressed upon them greatly; and they turned

in unto him, and entered into his house; and he made them a feast, and did bake unleavened bread, and they did eat.

los oppido ut deuerterent ad eum ingressisque domum illius fecit conuiuium coxit azyma et comederunt 19:4 prius 19:4 But before they lay down, the men of the city, even

the men of Sodom, compassed the house round, both old and young, all the people from every quarter:

autem quam irent cubitum uiri ciuitatis uallauerunt domum

a puero usque ad senem omnis populus simul 19:5 uoca- 19:5 And they called unto Lot, and said unto him, Where are the men which came in to thee this night? bring them out unto us, that we may know them.

ueruntque Loth et dixerunt ei ubi sunt uiri qui introierunt

ad te nocte educ illos huc ut cognoscamus eos 19:6 eg- 19:6 And Lot went out at the door unto them, and shut the

door after him,

ressus ad eos Loth post tergum adcludens ostium ait 19:7 19:7 And said, I pray you, brethren, do not so wickedly. nolite quaeso fratres mei nolite malum hoc facere 19:8 ha- 19:8 Behold now, I have two daughters which have not

known man; let me, I pray you, bring them out unto you, and do ye to them as is good in your eyes: only unto these men do nothing; for therefore came they under the shadow of my roof.

beo duas filias quae necdum cognouerunt uirum educam eas ad uos et abutimini eis sicut placuerit uobis dummodo uiris istis nihil faciatis mali quia ingressi sunt sub umbraculum tegminis mei 19:9 at illi dixerunt recede illuc et rursus in- 19:9 And they said, Stand back. And they said again, This

one fellow came in to sojourn, and he will needs be a judge: now will we deal worse with thee, than with them. And they pressed sore upon the man, even Lot, and came near to break the door.

gressus es inquiunt ut aduena numquid ut iudices te ergo ipsum magis quam hos adfligemus uimque faciebant Loth uehementissime iam prope erat ut refringerent fores 19:10 19:10 But the men put forth their hand, and pulled Lot into

the house to them, and shut to the door.

et ecce miserunt manum uiri et introduxerunt ad se Loth cluseruntque ostium 19:11 et eos qui erant foris percusserunt 19:11 And they smote the men that were at the door of the

house with blindness, both small and great: so that they wearied themselves to find the door.

caecitate a minimo usque ad maximum ita ut ostium inuenire non possent 19:12 dixerunt autem ad Loth habes hic 19:12 And the men said unto Lot, Hast thou here any be-

sides? son in law, and thy sons, and thy daughters, and whatsoever thou hast in the city, bring them out of this

tuorum quempiam generum aut filios aut filias omnes qui tui place:

sunt educ de urbe hac 19:13 delebimus enim locum istum 19:13 For we will destroy this place, because the cry of them is waxen great before the face of the LORD; and the LORD hath sent us to destroy it.

eo quod increuerit clamor eorum coram Domino qui misit nos ut perdamus illos 19:14 egressus itaque Loth locutus 19:14 And Lot went out, and spake unto his sons in law,

which married his daughters, and said, Up, get you out of this place; for the LORD will destroy this city. But he

est ad generos suos qui accepturi erant filias eius et dixit seemed as one that mocked unto his sons in law. surgite egredimini de loco isto quia delebit Dominus ciuitatem hanc et uisus est eis quasi ludens loqui 19:15 cumque 19:15 And when the morning arose, then the angels has-

tened Lot, saying, Arise, take thy wife, and thy two daughters, which are here; lest thou be consumed in the iniquity

esset mane cogebant eum angeli dicentes surge et tolle uxo- of the city. rem tuam et duas filias quas habes ne et tu pariter pereas in

Biblia Sacra Vulgata

19:16 And while he lingered, the men laid hold upon his scelere ciuitatis 19:16 dissimulante illo adprehenderunt ma-

hand, and upon the hand of his wife, and upon the hand of his two daughters; the LORD being merciful unto him: and they brought him forth, and set him without the city. num eius et manum uxoris ac duarum filiarum eius eo quod

19:17 And it came to pass, when they had brought them parceret Dominus illi 19:17 et eduxerunt eum posueruntque

forth abroad, that he said, Escape for thy life; look not behind thee, neither stay thou in all the plain; escape to

the mountain, lest thou be consumed. extra ciuitatem ibi locutus est ad eum salua animam tuam

noli respicere post tergum nec stes in omni circa regione sed

19:18 And Lot said unto them, Oh, not so, my LORD: in monte saluum te fac ne et tu simul pereas 19:18 dixitque 19:19 Behold now, thy servant hath found grace in thy Loth ad eos quaeso Domine mi 19:19 quia inuenit seruus

sight, and thou hast magnified thy mercy, which thou hast shewed unto me in saving my life; and I cannot escape to

the mountain, lest some evil take me, and I die: tuus gratiam coram te et magnificasti misericordiam tuam quam fecisti mecum ut saluares animam meam nec possum in monte saluari ne forte adprehendat me malum et moriar

19:20 Behold now, this city is near to flee unto, and it is a 19:20 est ciuitas haec iuxta ad quam possum fugere parua et

little one: Oh, let me escape thither, (is it not a little one?) and my soul shall live.

saluabor in ea numquid non modica est et uiuet anima mea

19:21 And he said unto him, See, I have accepted thee 19:21 dixitque ad eum ecce etiam in hoc suscepi preces tuas

concerning this thing also, that I will not overthrow this city, for the which thou hast spoken.

ut non subuertam urbem pro qua locutus es 19:22 festina et

19:22 Haste thee, escape thither; for I cannot do anything till thou be come thither. Therefore the name of the city was called Zoar.

saluare ibi quia non potero facere quicquam donec ingredia-

19:23 The sun was risen upon the earth when Lot entered ris illuc idcirco uocatum est nomen urbis illius Segor 19:23

into Zoar.

sol egressus est super terram et Loth ingressus est in Se-

19:24 Then the LORD rained upon Sodom and upon Go- gor 19:24 igitur Dominus pluit super Sodomam et Gomor-

morrah brimstone and fire from the LORD out of heaven; 19:25 And he overthrew those cities, and all the plain, and ram sulphur et ignem a Domino de caelo 19:25 et subuertit all the inhabitants of the cities, and that which grew upon the ground.

ciuitates has et omnem circa regionem uniuersos habitatores

19:26 But his wife looked back from behind him, and she urbium et cuncta terrae uirentia 19:26 respiciensque uxor

became a pillar of salt.

19:27 And Abraham gat up early in the morning to the eius post se uersa est in statuam salis 19:27 Abraham au-

place where he stood before the LORD:

19:28 And he looked toward Sodom and Gomorrah, and tem consurgens mane ubi steterat prius cum Domino 19:28

toward all the land of the plain, and beheld, and, lo, the smoke of the country went up as the smoke of a furnace.

intuitus est Sodomam et Gomorram et uniuersam terram regionis illius uiditque ascendentem fauillam de terra quasi

19:29 And it came to pass, when God destroyed the cities fornacis fumum 19:29 cum enim subuerteret Deus ciuita-

of the plain, that God remembered Abraham, and sent Lot out of the midst of the overthrow, when he overthrew the

cities in the which Lot dwelt. tes regionis illius recordatus est Abrahae et liberauit Loth 19:30 And Lot went up out of Zoar, and dwelt in the moun- de subuersione urbium in quibus habitauerat 19:30 ascen-

tain, and his two daughters with him; for he feared to dwell in Zoar: and he dwelt in a cave, he and his two daughters.

ditque Loth de Segor et mansit in monte duae quoque filiae eius cum eo timuerat enim manere in Segor et mansit in spe-

19:31 And the firstborn said unto the younger, Our father lunca ipse et duae filiae eius 19:31 dixitque maior ad miis old, and there is not a man in the earth to come in unto us after the manner of all the earth:

norem pater noster senex est et nullus uirorum remansit in terra qui possit ingredi ad nos iuxta morem uniuersae ter-

Liber Genesis

rae 19:32 ueni inebriemus eum uino dormiamusque cum eo 19:32 Come, let us make our father drink wine, and we

will lie with him, that we may preserve seed of our father.

ut seruare possimus ex patre nostro semen 19:33 dederunt 19:33 And they made their father drink wine that night:

and the firstborn went in, and lay with her father; and he perceived not when she lay down, nor when she arose.

itaque patri suo bibere uinum nocte illa et ingressa est maior dormiuitque cum patre at ille non sensit nec quando accubuit filia nec quando surrexit 19:34 altera quoque die dixit 19:34 And it came to pass on the morrow, that the firstborn

said unto the younger, Behold, I lay yesternight with my father: let us make him drink wine this night also; and go thou in, and lie with him, that we may preserve seed of our father.

maior ad minorem ecce dormiui heri cum patre meo demus ei bibere uinum etiam hac nocte et dormies cum eo ut saluemus semen de patre nostro 19:35 dederunt et illa nocte patri 19:35 And they made their father drink wine that night

also: and the younger arose, and lay with him; and he perceived not when she lay down, nor when she arose.

uinum ingressaque minor filia dormiuit cum eo et nec tunc quidem sensit quando concubuerit uel quando illa surrexerit 19:36 conceperunt ergo duae filiae Loth de patre suo 19:37 19:36 Thus were both the daughters of Lot with child by

their father.

19:37 And the first born bare a son, and called his name Moab: the same is the father of the Moabites unto this day.

peperitque maior filium et uocauit nomen eius Moab ipse est

pater Moabitarum usque in praesentem diem 19:38 minor 19:38 And the younger, she also bare a son, and called his name Benammi: the same is the father of the children of Ammon unto this day.

quoque peperit filium et uocauit nomen eius Ammon id est filius populi mei ipse est pater Ammanitarum usque hodie

20:1 profectus inde Abraham in terram australem habita- 20:1 And Abraham journeyed from thence toward the

south country, and dwelled between Kadesh and Shur, and sojourned in Gerar.

uit inter Cades et Sur et peregrinatus est in Geraris 20:2

20:2 And Abraham said of Sarah his wife, She is my sister: and Abimelech king of Gerar sent, and took Sarah.

dixitque de Sarra uxore sua soror mea est misit ergo Abimelech rex Gerarae et tulit eam 20:3 uenit autem Deus ad 20:3 But God came to Abimelech in a dream by night,

and said to him, Behold, thou art but a dead man, for the woman which thou hast taken; for she is a man’s wife.

Abimelech per somnium noctis et ait ei en morieris propter mulierem quam tulisti habet enim uirum 20:4 Abimelech 20:4 But Abimelech had not come near her: and he said,

LORD, wilt thou slay also a righteous nation?

uero non tetigerat eam et ait Domine num gentem ignorantem et iustam interficies 20:5 nonne ipse dixit mihi soror 20:5 Said he not unto me, She is my sister? and she, even

she herself said, He is my brother: in the integrity of my heart and innocency of my hands have I done this.

mea est et ipsa ait frater meus est in simplicitate cordis mei et munditia manuum mearum feci hoc 20:6 dixitque ad eum 20:6 And God said unto him in a dream, Yea, I know that

thou didst this in the integrity of thy heart; for I also withheld thee from sinning against me: therefore suffered I

Deus et ego scio quod simplici corde feceris et ideo custo- thee not to touch her. diui te ne peccares in me et non dimisi ut tangeres eam 20:7 20:7 Now therefore restore the man his wife; for he is a

prophet, and he shall pray for thee, and thou shalt live: and if thou restore her not, know thou that thou shalt surely die,

nunc igitur redde uxorem uiro suo quia propheta est et ora- thou, and all that are thine. bit pro te et uiues si autem nolueris reddere scito quod morte morieris tu et omnia quae tua sunt 20:8 statimque de nocte 20:8 Therefore Abimelech rose early in the morning, and

called all his servants, and told all these things in their ears: and the men were sore afraid.

consurgens Abimelech uocauit omnes seruos suos et locutus est uniuersa uerba haec in auribus eorum timueruntque omnes uiri ualde 20:9 uocauit autem Abimelech etiam Ab- 20:9 Then Abimelech called Abraham, and said unto him,

What hast thou done unto us? and what have I offended thee, that thou hast brought on me and on my kingdom a great sin? thou hast done deeds unto me that ought not to be done.

Biblia Sacra Vulgata

raham et dixit ei quid fecisti nobis quid peccauimus in te quia induxisti super me et super regnum meum peccatum

20:10 And Abimelech said unto Abraham, What sawest grande quae non debuisti facere fecisti nobis 20:10 rursus-

thou, that thou hast done this thing?

20:11 And Abraham said, Because I thought, Surely the que expostulans ait quid uidisti ut hoc faceres 20:11 res-

fear of God is not in this place; and they will slay me for my wife’s sake.

pondit Abraham cogitaui mecum dicens forsitan non est timor Dei in loco isto et interficient me propter uxorem meam

20:12 And yet indeed she is my sister; she is the daughter 20:12 alias autem et uere soror mea est filia patris mei et non

of my father, but not the daughter of my mother; and she became my wife.

filia matris meae et duxi eam uxorem 20:13 postquam au-

20:13 And it came to pass, when God caused me to wander from my father’s house, that I said unto her, This is thy kindness which thou shalt shew unto me; at every place whither we shall come, say of me, He is my brother.

tem eduxit me Deus de domo patris mei dixi ad eam hanc misericordiam facies mecum in omni loco ad quem ingredie-

20:14 And Abimelech took sheep, and oxen, and menser- mur dices quod frater tuus sim 20:14 tulit igitur Abimelech

vants, and womenservants, and gave them unto Abraham, and restored him Sarah his wife.

oues et boues et seruos et ancillas et dedit Abraham reddidit-

20:15 And Abimelech said, Behold, my land is before que illi Sarram uxorem suam 20:15 et ait terra coram uobis

thee: dwell where it pleaseth thee.

20:16 And unto Sarah he said, Behold, I have given thy est ubicumque tibi placuerit habita 20:16 Sarrae autem dixit

brother a thousand pieces of silver: behold, he is to thee a covering of the eyes, unto all that are with thee, and with

all other: thus she was reproved. ecce mille argenteos dedi fratri tuo hoc erit tibi in uelamen

oculorum ad omnes qui tecum sunt et quocumque perrexeris

20:17 So Abraham prayed unto God: and God healed mementoque te deprehensam 20:17 orante autem Abraham

Abimelech, and his wife, and his maidservants; and they bare children.

sanauit Deus Abimelech et uxorem ancillasque eius et pepe-

20:18 For the LORD had fast closed up all the wombs rerunt 20:18 concluserat enim Deus omnem uuluam domus

of the house of Abimelech, because of Sarah Abraham’s wife.

Abimelech propter Sarram uxorem Abraham

21:1 And the LORD visited Sarah as he had said, and the 21:1 uisitauit autem Dominus Sarram sicut promiserat et

LORD did unto Sarah as he had spoken.

21:2 For Sarah conceived, and bare Abraham a son in his impleuit quae locutus est 21:2 concepitque et peperit filium

old age, at the set time of which God had spoken to him.

21:3 And Abraham called the name of his son that was in senectute sua tempore quo praedixerat ei Deus 21:3 uo-

born unto him, whom Sarah bare to him, Isaac.

cauitque Abraham nomen filii sui quem genuit ei Sarra Isaac

21:4 And Abraham circumcised his son Isaac being eight 21:4 et circumcidit eum octauo die sicut praeceperat ei De-

days old, as God had commanded him.

21:5 And Abraham was an hundred years old, when his us 21:5 cum centum esset annorum hac quippe aetate pat-

son Isaac was born unto him.

21:6 And Sarah said, God hath made me to laugh, so that ris natus est Isaac 21:6 dixitque Sarra risum fecit mihi De-

all that hear will laugh with me.

21:7 And she said, Who would have said unto Abraham, us quicumque audierit conridebit mihi 21:7 rursumque ait

that Sarah should have given children suck? for I have born him a son in his old age.

quis auditurum crederet Abraham quod Sarra lactaret filium

21:8 And the child grew, and was weaned: and Abraham quem peperit ei iam seni 21:8 creuit igitur puer et ablacta-

made a great feast the same day that Isaac was weaned.

tus est fecitque Abraham grande conuiuium in die ablacta-

21:9 And Sarah saw the son of Hagar the Egyptian, which tionis eius 21:9 cumque uidisset Sarra filium Agar Aegy-

she had born unto Abraham, mocking.

21:10 Wherefore she said unto Abraham, Cast out this ptiae ludentem dixit ad Abraham 21:10 eice ancillam hanc

bondwoman and her son: for the son of this bondwoman shall not be heir with my son, even with Isaac.

Liber Genesis

et filium eius non enim erit heres filius ancillae cum filio

meo Isaac 21:11 dure accepit hoc Abraham pro filio suo 21:11 And the thing was very grievous in Abraham’s sight

because of his son.

21:12 cui dixit Deus non tibi uideatur asperum super puero 21:12 And God said unto Abraham, Let it not be grievous

in thy sight because of the lad, and because of thy bondwoman; in all that Sarah hath said unto thee, hearken unto

et super ancilla tua omnia quae dixerit tibi Sarra audi uo- her voice; for in Isaac shall thy seed be called. cem eius quia in Isaac uocabitur tibi semen 21:13 sed et 21:13 And also of the son of the bondwoman will I make

a nation, because he is thy seed.

filium ancillae faciam in gentem magnam quia semen tuum est 21:14 surrexit itaque Abraham mane et tollens panem 21:14 And Abraham rose up early in the morning, and

took bread, and a bottle of water, and gave it unto Hagar, putting it on her shoulder, and the child, and sent her away: and she departed, and wandered in the wilderness of Beersheba.

et utrem aquae inposuit scapulae eius tradiditque puerum et dimisit eam quae cum abisset errabat in solitudine Bersabee 21:15 cumque consumpta esset aqua in utre abiecit pue- 21:15 And the water was spent in the bottle, and she cast

the child under one of the shrubs.

rum subter unam arborum quae ibi erant 21:16 et abiit sedit- 21:16 And she went, and sat her down over against him a

good way off, as it were a bow shot: for she said, Let me not see the death of the child. And she sat over against

que e regione procul quantum potest arcus iacere dixit enim him, and lift up her voice, and wept. non uidebo morientem puerum et sedens contra leuauit uocem suam et fleuit 21:17 exaudiuit autem Deus uocem pueri 21:17 And God heard the voice of the lad; and the angel

of God called to Hagar out of heaven, and said unto her, What aileth thee, Hagar? fear not; for God hath heard the

uocauitque angelus Domini Agar de caelo dicens quid agis voice of the lad where he is. Agar noli timere exaudiuit enim Deus uocem pueri de loco

in quo est 21:18 surge tolle puerum et tene manum illius 21:18 Arise, lift up the lad, and hold him in thine hand; for

I will make him a great nation.

quia in gentem magnam faciam eum 21:19 aperuitque ocu- 21:19 And God opened her eyes, and she saw a well of

water; and she went, and filled the bottle with water, and gave the lad drink.

los eius Deus quae uidens puteum aquae abiit et impleuit utrem deditque puero bibere 21:20 et fuit cum eo qui creuit 21:20 And God was with the lad; and he grew, and dwelt

in the wilderness, and became an archer.

et moratus est in solitudine et factus est iuuenis sagittarius 21:21 habitauitque in deserto Pharan et accepit illi mater sua 21:21 And he dwelt in the wilderness of Paran: and his

mother took him a wife out of the land of Egypt.

uxorem de terra Aegypti 21:22 eodem tempore dixit Abi- 21:22 And it came to pass at that time, that Abimelech and Phichol the chief captain of his host spake unto Abraham, saying, God is with thee in all that thou doest:

melech et Fichol princeps exercitus eius ad Abraham Deus tecum est in uniuersis quae agis 21:23 iura ergo per Do- 21:23 Now therefore swear unto me here by God that thou

wilt not deal falsely with me, nor with my son, nor with my son’s son: but according to the kindness that I have done unto thee, thou shalt do unto me, and to the land wherein thou hast sojourned.

minum ne noceas mihi et posteris meis stirpique meae sed iuxta misericordiam quam feci tibi facies mihi et terrae in qua uersatus es aduena 21:24 dixitque Abraham ego iurabo 21:24 And Abraham said, I will swear. 21:25 et increpauit Abimelech propter puteum aquae quem ui 21:25 And Abraham reproved Abimelech because of a

well of water, which Abimelech’s servants had violently taken away.

abstulerant serui illius 21:26 respondit Abimelech nesciui

21:26 And Abimelech said, I wot not who hath done this thing; neither didst thou tell me, neither yet heard I of it, but to day.

quis fecerit hanc rem sed et tu non indicasti mihi et ego non audiui praeter hodie 21:27 tulit itaque Abraham oues et bo- 21:27 And Abraham took sheep and oxen, and gave them

unto Abimelech; and both of them made a covenant.

ues et dedit Abimelech percusseruntque ambo foedus 21:28 21:28 And Abraham set seven ewe lambs of the flock by

themselves.

Biblia Sacra Vulgata

21:29 And Abimelech said unto Abraham, What mean et statuit Abraham septem agnas gregis seorsum 21:29 cui

these seven ewe lambs which thou hast set by themselves?

dixit Abimelech quid sibi uolunt septem agnae istae quas

21:30 And he said, For these seven ewe lambs shalt thou stare fecisti seorsum 21:30 at ille septem inquit agnas ac-

take of my hand, that they may be a witness unto me, that I have digged this well.

cipies de manu mea ut sint in testimonium mihi quoniam

21:31 Wherefore he called that place Beersheba; because ego fodi puteum istum 21:31 idcirco uocatus est locus ille

there they sware both of them.

21:32 Thus they made a covenant at Beersheba: then Bersabee quia ibi uterque iurauerunt 21:32 et inierunt foe-

Abimelech rose up, and Phichol the chief captain of his host, and they returned into the land of the Philistines.

dus pro puteo Iuramenti 21:33 surrexit autem Abimelech et

21:33 And Abraham planted a grove in Beersheba, and called there on the name of the LORD, the everlasting God.

Fichol princeps militiae eius reuersique sunt in terram Palestinorum Abraham uero plantauit nemus in Bersabee et in-

21:34 And Abraham sojourned in the Philistines’ land uocauit ibi nomen Domini Dei aeterni 21:34 et fuit colonus

many days.

terrae Philisthinorum diebus multis

22:1 And it came to pass after these things, that God did 22:1 quae postquam gesta sunt temptauit Deus Abraham et

tempt Abraham, and said unto him, Abraham: and he said, Behold, here I am.

dixit ad eum Abraham ille respondit adsum 22:2 ait ei tolle

22:2 And he said, Take now thy son, thine only son Isaac, whom thou lovest, and get thee into the land of Moriah; and offer him there for a burnt offering upon one of the mountains which I will tell thee of.

filium tuum unigenitum quem diligis Isaac et uade in terram Visionis atque offer eum ibi holocaustum super unum mon-

22:3 And Abraham rose up early in the morning, and sad- tium quem monstrauero tibi 22:3 igitur Abraham de nocte

dled his ass, and took two of his young men with him, and Isaac his son, and clave the wood for the burnt offering, and rose up, and went unto the place of which God had told him.

consurgens strauit asinum suum ducens secum duos iuuenes et Isaac filium suum cumque concidisset ligna in holocaus-

22:4 Then on the third day Abraham lifted up his eyes, and tum abiit ad locum quem praeceperat ei Deus 22:4 die ausaw the place afar off.

22:5 And Abraham said unto his young men, Abide ye tem tertio eleuatis oculis uidit locum procul 22:5 dixitque

here with the ass; and I and the lad will go yonder and worship, and come again to you.

ad pueros suos expectate hic cum asino ego et puer illuc usque properantes postquam adorauerimus reuertemur ad uos

22:6 And Abraham took the wood of the burnt offering, 22:6 tulit quoque ligna holocausti et inposuit super Isaac fi-

and laid it upon Isaac his son; and he took the fire in his hand, and a knife; and they went both of them together.

lium suum ipse uero portabat in manibus ignem et gladium

22:7 And Isaac spake unto Abraham his father, and said, cumque duo pergerent simul 22:7 dixit Isaac patri suo pater

My father: and he said, Here am I, my son. And he said, Behold the fire and the wood: but where is the lamb for a

burnt offering? mi at ille respondit quid uis fili ecce inquit ignis et ligna ubi 22:8 And Abraham said, My son, God will provide him- est uictima holocausti 22:8 dixit Abraham Deus prouidebit

self a lamb for a burnt offering: so they went both of them together.

sibi uictimam holocausti fili mi pergebant ergo pariter 22:9

22:9 And they came to the place which God had told him of; and Abraham built an altar there, and laid the wood in order, and bound Isaac his son, and laid him on the altar upon the wood.

ueneruntque ad locum quem ostenderat ei Deus in quo aedificauit altare et desuper ligna conposuit cumque conligasset Isaac filium suum posuit eum in altari super struem lignorum

22:10 And Abraham stretched forth his hand, and took the 22:10 extenditque manum et arripuit gladium ut immolaret fi-

knife to slay his son.

22:11 And the angel of the LORD called unto him out of lium 22:11 et ecce angelus Domini de caelo clamauit dicens

heaven, and said, Abraham, Abraham: and he said, Here am I.

Liber Genesis

Abraham Abraham qui respondit adsum 22:12 dixitque ei 22:12 And he said, Lay not thine hand upon the lad, nei-

ther do thou any thing unto him: for now I know that thou fearest God, seeing thou hast not withheld thy son, thine

non extendas manum tuam super puerum neque facias illi only son from me. quicquam nunc cognoui quod timeas Dominum et non peperceris filio tuo unigenito propter me 22:13 leuauit Abra- 22:13 And Abraham lifted up his eyes, and looked, and

behold behind him a ram caught in a thicket by his horns: and Abraham went and took the ram, and offered him up

ham oculos uiditque post tergum arietem inter uepres heren- for a burnt offering in the stead of his son. tem cornibus quem adsumens obtulit holocaustum pro filio 22:14 appellauitque nomen loci illius Dominus uidet unde us- 22:14 And Abraham called the name of that place Jeho-

vahjireh: as it is said to this day, In the mount of the LORD it shall be seen.

que hodie dicitur in monte Dominus uidebit 22:15 uocauit

22:15 And the angel of the LORD called unto Abraham out of heaven the second time,

autem angelus Domini Abraham secundo de caelo dicens 22:16 per memet ipsum iuraui dicit Dominus quia fecisti rem 22:16 And said, By myself have I sworn, saith the LORD,

for because thou hast done this thing, and hast not withheld thy son, thine only son:

hanc et non pepercisti filio tuo unigenito 22:17 benedicam

22:17 That in blessing I will bless thee, and in multiplying I will multiply thy seed as the stars of the heaven, and as the sand which is upon the sea shore; and thy seed shall possess the gate of his enemies;

tibi et multiplicabo semen tuum sicut stellas caeli et uelut harenam quae est in litore maris possidebit semen tuum portas inimicorum suorum 22:18 et benedicentur in semine tuo 22:18 And in thy seed shall all the nations of the earth be

blessed; because thou hast obeyed my voice.

omnes gentes terrae quia oboedisti uoci meae 22:19 reuer- 22:19 So Abraham returned unto his young men, and they

rose up and went together to Beersheba; and Abraham dwelt at Beersheba.

sus est Abraham ad pueros suos abieruntque Bersabee simul et habitauit ibi 22:20 his itaque gestis nuntiatum est Abra- 22:20 And it came to pass after these things, that it was

told Abraham, saying, Behold, Milcah, she hath also born children unto thy brother Nahor;

ham quod Melcha quoque genuisset filios Nahor fratri suo 22:21 Hus primogenitum et Buz fratrem eius Camuhel pat- 22:21 Huz his firstborn, and Buz his brother, and Kemuel

the father of Aram,

rem Syrorum 22:22 et Chased et Azau Pheldas quoque et 22:22 And Chesed, and Hazo, and Pildash, and Jidlaph,

and Bethuel.

Iedlaph 22:23 ac Bathuel de quo nata est Rebecca octo is- 22:23 And Bethuel begat Rebekah: these eight Milcah did

bear to Nahor, Abraham’s brother.

tos genuit Melcha Nahor fratri Abraham 22:24 concubina 22:24 And his concubine, whose name was Reumah, she

bare also Tebah, and Gaham, and Thahash, and Maachah.

uero illius nomine Roma peperit Tabee et Gaom et Thaas et Maacha

23:1 uixit autem Sarra centum uiginti septem annis 23:2 23:1 And Sarah was an hundred and seven and twenty

years old: these were the years of the life of Sarah.

23:2 And Sarah died in Kirjatharba; the same is Hebron in the land of Canaan: and Abraham came to mourn for Sarah, and to weep for her.

et mortua est in ciuitate Arbee quae est Hebron in terra Chanaan uenitque Abraham ut plangeret et fleret eam 23:3 _{23:3 And Abraham stood up from before his dead, and}

spake unto the sons of Heth, saying,

cumque surrexisset ab officio funeris locutus est ad filios Heth dicens 23:4 aduena sum et peregrinus apud uos date 23:4 I am a stranger and a sojourner with you: give me a

possession of a buryingplace with you, that I may bury my dead out of my sight.

mihi ius sepulchri uobiscum ut sepeliam mortuum meum 23:5 responderuntque filii Heth 23:6 audi nos domine prin- 23:5 And the children of Heth answered Abraham, saying

unto him,

23:6 Hear us, my lord: thou art a mighty prince among us: in the choice of our sepulchres bury thy dead; none of us shall withhold from thee his sepulchre, but that thou mayest bury thy dead.

ceps Dei es apud nos in electis sepulchris nostris sepeli mortuum tuum nullusque prohibere te poterit quin in mo-

Biblia Sacra Vulgata

23:7 And Abraham stood up, and bowed himself to the numento eius sepelias mortuum tuum 23:7 surrexit Abra-

people of the land, even to the children of Heth.

23:8 And he communed with them, saying, If it be your ham et adorauit populum terrae filios uidelicet Heth 23:8

mind that I should bury my dead out of my sight; hear me, and intreat for me to Ephron the son of Zohar,

dixitque ad eos si placet animae uestrae ut sepeliam mortuum meum audite me et intercedite apud Ephron filium

23:9 That he may give me the cave of Machpelah, which Soor 23:9 ut det mihi speluncam duplicem quam habet in

he hath, which is in the end of his field; for as much money as it is worth he shall give it me for a possession of a bury-

ingplace amongst you. extrema parte agri sui pecunia digna tradat mihi eam co-23:10 And Ephron dwelt among the children of Heth: and ram uobis in possessionem sepulchri 23:10 habitabat autem

Ephron the Hittite answered Abraham in the audience of the children of Heth, even of all that went in at the gate of

his city, saying, Ephron in medio filiorum Heth responditque ad Abraham

cunctis audientibus qui ingrediebantur portam ciuitatis illius

23:11 Nay, my lord, hear me: the field give I thee, and the dicens 23:11 nequaquam ita fiat domine mi sed magis aucave that is therein, I give it thee; in the presence of the sons of my people give I it thee: bury thy dead.

sculta quod loquor agrum trado tibi et speluncam quae in eo est praesentibus filiis populi mei sepeli mortuum tuum

23:12 And Abraham bowed down himself before the peo- 23:12 adorauit Abraham coram populo terrae 23:13 et locu-

ple of the land. 23:13 And he spake unto Ephron in the audience of the people of the land, saying, But if thou wilt give it, I pray thee, hear me: I will give thee money for the field; take it of me, and I will bury my dead there.

tus est ad Ephron circumstante plebe quaeso ut audias me dabo pecuniam pro agro suscipe eam et sic sepeliam mor-

23:14 And Ephron answered Abraham, saying unto him, tuum meum in eo 23:14 respondit Ephron 23:15 domine 23:15 My lord, hearken unto me: the land is worth four hundred shekels of silver; what is that betwixt me and thee? bury therefore thy dead.

mi audi terram quam postulas quadringentis argenti siclis ualet istud est pretium inter me et te sed quantum est hoc

23:16 And Abraham hearkened unto Ephron; and Abra- sepeli mortuum tuum 23:16 quod cum audisset Abraham

ham weighed to Ephron the silver, which he had named in the audience of the sons of Heth, four hundred shekels of

silver, current money with the merchant. adpendit pecuniam quam Ephron postulauerat audientibus

filiis Heth quadringentos siclos argenti et probati monetae

23:17 And the field of Ephron which was in Machpelah, publicae 23:17 confirmatusque est ager quondam Ephro-

which was before Mamre, the field, and the cave which was therein, and all the trees that were in the field, that

were in all the borders round about, were made sure nis in quo erat spelunca duplex respiciens Mambre tam ipse

quam spelunca et omnes arbores eius in cunctis terminis per

23:18 Unto Abraham for a possession in the presence of circuitum 23:18 Abrahae in possessionem uidentibus filiis

the children of Heth, before all that went in at the gate of his city.

Heth et cunctis qui intrabant portam ciuitatis illius 23:19 at-

23:19 And after this, Abraham buried Sarah his wife in the cave of the field of Machpelah before Mamre: the same is Hebron in the land of Canaan.

que ita sepeliuit Abraham Sarram uxorem suam in spelunca agri duplici qui respiciebat Mambre haec est Hebron in terra

23:20 And the field, and the cave that is therein, were Chanaan 23:20 et confirmatus est ager et antrum quod erat

made sure unto Abraham for a possession of a buryingplace by the sons of Heth.

in eo Abrahae in possessionem monumenti a filiis Heth

24:1 And Abraham was old, and well stricken in age: and 24:1 erat autem Abraham senex dierumque multorum et the LORD had blessed Abraham in all things.

24:2 And Abraham said unto his eldest servant of his Dominus in cunctis benedixerat ei 24:2 dixitque ad seruum

house, that ruled over all that he had, Put, I pray thee, thy hand under my thigh:

seniorem domus suae qui praeerat omnibus quae habebat

Liber Genesis

pone manum tuam subter femur meum 24:3 ut adiurem te 24:3 And I will make thee swear by the LORD, the God of

heaven, and the God of the earth, that thou shalt not take a wife unto my son of the daughters of the Canaanites,

per Dominum Deum caeli et terrae ut non accipias uxorem among whom I dwell:

filio meo de filiabus Chananeorum inter quos habito 24:4 24:4 But thou shalt go unto my country, and to my kindred,

and take a wife unto my son Isaac.

sed ad terram et ad cognationem meam proficiscaris et inde accipias uxorem filio meo Isaac 24:5 respondit seruus si 24:5 And the servant said unto him, Peradventure the

woman will not be willing to follow me unto this land: must I needs bring thy son again unto the land from

noluerit mulier uenire mecum in terram hanc num reducere whence thou camest? debeo filium tuum ad locum de quo egressus es 24:6 dixit 24:6 And Abraham said unto him, Beware thou that thou

bring not my son thither again.

Abraham caue nequando reducas illuc filium meum 24:7 24:7 The LORD God of heaven, which took me from

my father’s house, and from the land of my kindred, and which spake unto me, and that sware unto me, saying, Unto thy seed will I give this land; he shall send his angel before thee, and thou shalt take a wife unto my son from thence.

Dominus Deus caeli qui tulit me de domo patris mei et de terra natiuitatis meae qui locutus est mihi et iurauit dicens semini tuo dabo terram hanc ipse mittet angelum suum coram te et accipies inde uxorem filio meo 24:8 sin autem no- 24:8 And if the woman will not be willing to follow thee,

then thou shalt be clear from this my oath: only bring not my son thither again.

luerit mulier sequi te non teneberis iuramento filium tantum

meum ne reducas illuc 24:9 posuit ergo seruus manum sub 24:9 And the servant put his hand under the thigh of Abra-

ham his master, and sware to him concerning that matter.

femore Abraham domini sui et iurauit illi super sermone hoc 24:10 tulitque decem camelos de grege domini sui et abiit 24:10 And the servant took ten camels of the camels of his

master, and departed; for all the goods of his master were in his hand: and he arose, and went to Mesopotamia, unto

ex omnibus bonis eius portans secum profectusque perrexit the city of Nahor. Mesopotamiam ad urbem Nahor 24:11 cumque camelos fe- 24:11 And he made his camels to kneel down without the

city by a well of water at the time of the evening, even the time that women go out to draw water.

cisset accumbere extra oppidum iuxta puteum aquae uespere eo tempore quo solent mulieres egredi ad hauriendam aquam dixit 24:12 Domine Deus domini mei Abraham occurre ob- 24:12 And he said O LORD God of my master Abraham, I

pray thee, send me good speed this day, and shew kindness unto my master Abraham.

secro hodie mihi et fac misericordiam cum domino meo Abraham 24:13 ecce ego sto propter fontem aquae et filiae ha- 24:13 Behold, I stand here by the well of water; and the

daughters of the men of the city come out to draw water:

bitatorum huius ciuitatis egredientur ad hauriendam aquam 24:14 igitur puella cui ego dixero inclina hydriam tuam ut bi- 24:14 And let it come to pass, that the damsel to whom

I shall say, Let down thy pitcher, I pray thee, that I may drink; and she shall say, Drink, and I will give thy camels drink also: let the same be she that thou hast appointed for thy servant Isaac; and thereby shall I know that thou hast shewed kindness unto my master.

bam et illa responderit bibe quin et camelis tuis dabo potum ipsa est quam praeparasti seruo tuo Isaac et per hoc intellegam quod feceris misericordiam cum domino meo 24:15 24:15 And it came to pass, before he had done speaking,

that, behold, Rebekah came out, who was born to Bethuel, son of Milcah, the wife of Nahor, Abraham’s brother, with

necdum intra se uerba conpleuerat et ecce Rebecca egredie- her pitcher upon her shoulder. batur filia Bathuel filii Melchae uxoris Nahor fratris Abraham habens hydriam in scapula 24:16 puella decora nimis 24:16 And the damsel was very fair to look upon, a virgin,

neither had any man known her: and she went down to the well, and filled her pitcher, and came up.

uirgoque pulcherrima et incognita uiro descenderat autem ad fontem et impleuerat hydriam ac reuertebatur 24:17 occur- 24:17 And the servant ran to meet her, and said, Let me, I

pray thee, drink a little water of thy pitcher.

Biblia Sacra Vulgata

ritque ei seruus et ait pauxillum mihi ad sorbendum prae-

24:18 And she said, Drink, my lord: and she hasted, and be aquae de hydria tua 24:18 quae respondit bibe domine

let down her pitcher upon her hand, and gave him drink.

mi celeriterque deposuit hydriam super ulnam suam et de-

24:19 And when she had done giving him drink, she said, dit ei potum 24:19 cumque ille bibisset adiecit quin et ca-

I will draw water for thy camels also, until they have done drinking.

melis tuis hauriam aquam donec cuncti bibant 24:20 effun-

24:20 And she hasted, and emptied her pitcher into the trough, and ran again unto the well to draw water, and drew for all his camels.

densque hydriam in canalibus recurrit ad puteum ut hauriret

24:21 And the man wondering at her held his peace, to wit aquam et haustam omnibus camelis dedit 24:21 ille autem

whether the LORD had made his journey prosperous or not.

contemplabatur eam tacitus scire uolens utrum prosperum

24:22 And it came to pass, as the camels had done drink- fecisset iter suum Dominus an non 24:22 postquam ergo ing, that the man took a golden earring of half a shekel weight, and two bracelets for her hands of ten shekels

weight of gold; biberunt cameli protulit uir inaures aureas adpendentes si-24:23 And said, Whose daughter art thou? tell me, I pray clos duos et armillas totidem pondo siclorum decem 24:23

thee: is there room in thy father’s house for us to lodge in?

dixitque ad eam cuius es filia indica mihi est in domo patris

24:24 And she said unto him, I am the daughter of Bethuel tui locus ad manendum 24:24 quae respondit filia Bathue-

the son of Milcah, which she bare unto Nahor.

24:25 She said moreover unto him, We have both straw lis sum filii Melchae quem peperit Nahor 24:25 et addidit

and provender enough, and room to lodge in.

dicens palearum quoque et faeni plurimum est apud nos et

24:26 And the man bowed down his head, and worshipped locus spatiosus ad manendum 24:26 inclinauit se homo et

the LORD.

24:27 And he said, Blessed be the LORD God of my mas- adorauit Dominum 24:27 dicens benedictus Dominus Deus

ter Abraham, who hath not left destitute my master of his mercy and his truth: I being in the way, the LORD led me

to the house of my master’s brethren. domini mei Abraham qui non abstulit misericordiam et ue-

ritatem suam a domino meo et recto me itinere perduxit in

24:28 And the damsel ran, and told them of her mother’s domum fratris domini mei 24:28 cucurrit itaque puella et

house these things.

24:29 And Rebekah had a brother, and his name was La- nuntiauit in domum matris suae omnia quae audierat 24:29

ban: and Laban ran out unto the man, unto the well.

habebat autem Rebecca fratrem nomine Laban qui festinus

24:30 And it came to pass, when he saw the earring and egressus est ad hominem ubi erat fons 24:30 cumque ui-

bracelets upon his sister’s hands, and when he heard the words of Rebekah his sister, saying, Thus spake the man unto me; that he came unto the man; and, behold, he stood by the camels at the well.

disset inaures et armillas in manibus sororis suae et audisset cuncta uerba referentis haec locutus est mihi homo uenit ad uirum qui stabat iuxta camelos et propter fontem aquae

24:31 And he said, Come in, thou blessed of the LORD; 24:31 dixitque ad eum ingredere benedicte Domini cur foris

wherefore standest thou without? for I have prepared the house, and room for the camels.

stas praeparaui domum et locum camelis 24:32 et introduxit

24:32 And the man came into the house: and he ungirded his camels, and gave straw and provender for the camels, and water to wash his feet, and the men’s feet that were with him.

eum hospitium ac destrauit camelos deditque paleas et faenum et aquam ad lauandos pedes camelorum et uirorum qui

24:33 And there was set meat before him to eat: but he uenerant cum eo 24:33 et adpositus est in conspectu eius

said, I will not eat, until I have told mine errand. And he said, Speak on.

panis qui ait non comedam donec loquar sermones meos re-

24:34 And he said, I am Abraham’s servant. spondit ei loquere 24:34 at ille seruus inquit Abraham sum

Liber Genesis

24:35 et Dominus benedixit domino meo ualde magnificatus- 24:35 And the LORD hath blessed my master greatly; and

he is become great: and he hath given him flocks, and herds, and silver, and gold, and menservants, and maidser-

que est et dedit ei oues et boues argentum et aurum seruos vants, and camels, and asses. et ancillas camelos et asinos 24:36 et peperit Sarra uxor do- 24:36 And Sarah my master’s wife bare a son to my master

when she was old: and unto him hath he given all that he hath.

mini mei filium domino meo in senectute sua deditque illi omnia quae habuerat 24:37 et adiurauit me dominus meus 24:37 And my master made me swear, saying, Thou shalt

not take a wife to my son of the daughters of the Canaanites, in whose land I dwell:

dicens non accipies uxorem filio meo de filiabus Chanane-

orum in quorum terra habito 24:38 sed ad domum patris 24:38 But thou shalt go unto my father’s house, and to my

kindred, and take a wife unto my son.

mei perges et de cognatione mea accipies uxorem filio meo 24:39 ego uero respondi domino meo quid si noluerit uenire 24:39 And I said unto my master, Peradventure the woman

will not follow me.

mecum mulier 24:40 Dominus ait in cuius conspectu am- 24:40 And he said unto me, The LORD, before whom I

walk, will send his angel with thee, and prosper thy way; and thou shalt take a wife for my son of my kindred, and

bulo mittet angelum suum tecum et diriget uiam tuam ac- of my father’s house: cipiesque uxorem filio meo de cognatione mea et de domo patris mei 24:41 innocens eris a maledictione mea cum ue- 24:41 Then shalt thou be clear from this my oath, when

thou comest to my kindred; and if they give not thee one, thou shalt be clear from my oath.

neris ad propinquos meos et non dederint tibi 24:42 ueni

24:42 And I came this day unto the well, and said, O LORD God of my master Abraham, if now thou do prosper my way which I go:

ergo hodie ad fontem et dixi Domine Deus domini mei Ab-

raham si direxisti uiam meam in qua nunc ambulo 24:43 24:43 Behold, I stand by the well of water; and it shall come to pass, that when the virgin cometh forth to draw water, and I say to her, Give me, I pray thee, a little water

ecce sto iuxta fontem aquae et uirgo quae egredietur ad hau- of thy pitcher to drink; riendam aquam audierit a me da mihi pauxillum aquae ad bibendum ex hydria tua 24:44 et dixerit mihi et tu bibe et 24:44 And she say to me, Both drink thou, and I will also

draw for thy camels: let the same be the woman whom the LORD hath appointed out for my master’s son.

camelis tuis hauriam ipsa est mulier quam praeparauit Dominus filio domini mei 24:45 dum haec mecum tacitus uol- 24:45 And before I had done speaking in mine heart, be-

hold, Rebekah came forth with her pitcher on her shoulder; and she went down unto the well, and drew water:

uerem apparuit Rebecca ueniens cum hydria quam portabat and I said unto her, Let me drink, I pray thee. in scapula descenditque ad fontem et hausit aquam et aio ad eam da mihi paululum bibere 24:46 quae festina depo- 24:46 And she made haste, and let down her pitcher from

her shoulder, and said, Drink, and I will give thy camels drink also: so I drank, and she made the camels drink also.

suit hydriam de umero et dixit mihi et tu bibe et camelis tuis potum tribuam bibi et adaquauit camelos 24:47 interroga- 24:47 And I asked her, and said, Whose daughter art thou?

And she said, the daughter of Bethuel, Nahor’s son, whom Milcah bare unto him: and I put the earring upon her face,

uique eam et dixi cuius es filia quae respondit filia Bathue- and the bracelets upon her hands. lis sum filii Nahor quem peperit illi Melcha suspendi itaque inaures ad ornandam faciem eius et armillas posui in manibus 24:48 pronusque adoraui Dominum benedicens Do- 24:48 And I bowed down my head, and worshipped the

LORD, and blessed the LORD God of my master Abraham, which had led me in the right way to take my mas-

mino Deo domini mei Abraham qui perduxisset me recto ter’s brother’s daughter unto his son. itinere ut sumerem filiam fratris domini mei filio eius 24:49 24:49 And now if ye will deal kindly and truly with my

master, tell me: and if not, tell me; that I may turn to the right hand, or to the left.

quam ob rem si facitis misericordiam et ueritatem cum do-

Biblia Sacra Vulgata

mino meo indicate mihi sin autem aliud placet et hoc dicite

24:50 Then Laban and Bethuel answered and said, The ut uadam ad dextram siue ad sinistram 24:50 responderunt

thing proceedeth from the LORD: we cannot speak unto thee bad or good.

Laban et Bathuel a Domino egressus est sermo non possu-

24:51 Behold, Rebekah is before thee, take her, and go, mus extra placitum eius quicquam aliud tecum loqui 24:51

and let her be thy master’s son’s wife, as the LORD hath spoken.

en Rebecca coram te est tolle eam et proficiscere et sit uxor

24:52 And it came to pass, that, when Abraham’s servant filii domini tui sicut locutus est Dominus 24:52 quod cum

heard their words, he worshipped the LORD, bowing himself to the earth.

audisset puer Abraham adorauit in terra Dominum 24:53

24:53 And the servant brought forth jewels of silver, and jewels of gold, and raiment, and gave them to Rebekah: he gave also to her brother and to her mother precious things.

prolatisque uasis argenteis et aureis ac uestibus dedit ea Rebeccae pro munere fratribus quoque eius et matri dona ob-

24:54 And they did eat and drink, he and the men that tulit 24:54 initoque conuiuio uescentes pariter et bibentes

were with him, and tarried all night; and they rose up in the morning, and he said, Send me away unto my master.

manserunt ibi surgens autem mane locutus est puer dimit-

24:55 And her brother and her mother said, Let the damsel tite me ut uadam ad dominum meum 24:55 responderunt abide with us a few days, at the least ten; after that she shall go.

fratres eius et mater maneat puella saltem decem dies apud

24:56 And he said unto them, Hinder me not, seeing the nos et postea proficiscetur 24:56 nolite ait me retinere quia

LORD hath prospered my way; send me away that I may go to my master.

Dominus direxit uiam meam dimittite me ut pergam ad do-

24:57 And they said, We will call the damsel, and enquire minum meum 24:57 dixerunt uocemus puellam et quaeraat her mouth.

24:58 And they called Rebekah, and said unto her, Wilt mus ipsius uoluntatem 24:58 cumque uocata uenisset sci-

thou go with this man? And she said, I will go.

24:59 And they sent away Rebekah their sister, and her scitati sunt uis ire cum homine isto quae ait uadam 24:59

nurse, and Abraham’s servant, and his men.

dimiserunt ergo eam et nutricem illius seruumque Abraham

24:60 And they blessed Rebekah, and said unto her, Thou et comites eius 24:60 inprecantes prospera sorori suae at-

art our sister, be thou the mother of thousands of millions, and let thy seed possess the gate of those which hate them.

que dicentes soror nostra es crescas in mille milia et pos-

24:61 And Rebekah arose, and her damsels, and they rode sideat semen tuum portas inimicorum suorum 24:61 igitur

upon the camels, and followed the man: and the servant took Rebekah, and went his way.

Rebecca et puellae illius ascensis camelis secutae sunt uirum

24:62 And Isaac came from the way of the well Lahairoi; qui festinus reuertebatur ad dominum suum 24:62 eo tem-

for he dwelt in the south country.

pore Isaac deambulabat per uiam quae ducit ad puteum cuius nomen est Viuentis et uidentis habitabat enim in terra au-

24:63 And Isaac went out to meditate in the field at the strali 24:63 et egressus fuerat ad meditandum in agro incli-

eventide: and he lifted up his eyes, and saw, and, behold, the camels were coming.

nata iam die cumque leuasset oculos uidit camelos uenientes

24:64 And Rebekah lifted up her eyes, and when she saw procul 24:64 Rebecca quoque conspecto Isaac descendit de

Isaac, she lighted off the camel. 24:65 For she had said unto the servant, What man is this camelo 24:65 et ait ad puerum quis est ille homo qui uethat walketh in the field to meet us? And the servant had said, It is my master: therefore she took a vail, and covered

herself. nit per agrum in occursum nobis dixit ei ipse est dominus 24:66 And the servant told Isaac all things that he had meus at illa tollens cito pallium operuit se 24:66 seruus au-

done.

24:67 And Isaac brought her into his mother Sarah’s tent, tem cuncta quae gesserat narrauit Isaac 24:67 qui introduxit

and took Rebekah, and she became his wife; and he loved her: and Isaac was comforted after his mother’s death.

Liber Genesis

eam in tabernaculum Sarrae matris suae et accepit uxorem et in tantum dilexit ut dolorem qui ex morte matris acciderat temperaret

25:1 Abraham uero aliam duxit uxorem nomine Cetthuram 25:1 Then again Abraham took a wife, and her name was

Keturah.

25:2 quae peperit ei Zamram et Iexan et Madan et Madian et 25:2 And she bare him Zimran, and Jokshan, and Medan,

and Midian, and Ishbak, and Shuah.

Iesboch et Sue 25:3 Iexan quoque genuit Saba et Dadan fi- 25:3 And Jokshan begat Sheba, and Dedan. And the sons

of Dedan were Asshurim, and Letushim, and Leummim.

lii Dadan fuerunt Assurim et Lathusim et Loommim 25:4 25:4 And the sons of Midian; Ephah, and Epher, and

Hanoch, and Abidah, and Eldaah. All these were the children of Keturah.

at uero ex Madian ortus est Epha et Opher et Enoch et Abida et Eldaa omnes hii filii Cetthurae 25:5 deditque Abraham 25:5 And Abraham gave all that he had unto Isaac. cuncta quae possederat Isaac 25:6 filiis autem concubina- 25:6 But unto the sons of the concubines, which Abraham

had, Abraham gave gifts, and sent them away from Isaac his son, while he yet lived, eastward, unto the east country.

rum largitus est munera et separauit eos ab Isaac filio suo dum adhuc ipse uiueret ad plagam orientalem 25:7 fuerunt 25:7 And these are the days of the years of Abraham’s life

which he lived, an hundred threescore and fifteen years.

autem dies uitae eius centum septuaginta quinque anni 25:8 25:8 Then Abraham gave up the ghost, and died in a good

old age, an old man, and full of years; and was gathered to his people.

et deficiens mortuus est in senectute bona prouectaeque aetatis et plenus dierum congregatusque est ad populum suum 25:9 et sepelierunt eum Isaac et Ismahel filii sui in spelunca 25:9 And his sons Isaac and Ishmael buried him in the cave

of Machpelah, in the field of Ephron the son of Zohar the Hittite, which is before Mamre;

duplici quae sita est in agro Ephron filii Soor Hetthei e regione Mambre 25:10 quem emerat a filiis Heth ibi sepultus 25:10 The field which Abraham purchased of the sons of

Heth: there was Abraham buried, and Sarah his wife.

est ipse et Sarra uxor eius 25:11 et post obitum illius bene- 25:11 And it came to pass after the death of Abraham, that

God blessed his son Isaac; and Isaac dwelt by the well Lahairoi.

dixit Deus Isaac filio eius qui habitabat iuxta puteum nomine Viuentis et uidentis 25:12 hae sunt generationes Ismahel fi- 25:12 Now these are the generations of Ishmael, Abra-

ham’s son, whom Hagar the Egyptian, Sarah’s handmaid, bare unto Abraham:

lii Abraham quem peperit ei Agar Aegyptia famula Sarrae 25:13 et haec nomina filiorum eius in uocabulis et genera- 25:13 And these are the names of the sons of Ishmael, by

their names, according to their generations: the firstborn of Ishmael, Nebajoth; and Kedar, and Adbeel, and Mib-

tionibus suis primogenitus Ismahelis Nabaioth dein Cedar et sam, Abdeel et Mabsam 25:14 Masma quoque et Duma et Massa 25:14 And Mishma, and Dumah, and Massa, 25:15 Adad et Thema Itur et Naphis et Cedma 25:16 isti sunt 25:15 Hadar, and Tema, Jetur, Naphish, and Kedemah:

25:16 These are the sons of Ishmael, and these are their names, by their towns, and by their castles; twelve princes according to their nations.

filii Ismahel et haec nomina per castella et oppida eorum duodecim principes tribuum suarum 25:17 anni uitae Isma- 25:17 And these are the years of the life of Ishmael, an

hundred and thirty and seven years: and he gave up the ghost and died; and was gathered unto his people.

hel centum triginta septem deficiens mortuus est et adpositus ad populum suum 25:18 habitauit autem ab Euila usque Sur 25:18 And they dwelt from Havilah unto Shur, that is be-

fore Egypt, as thou goest toward Assyria: and he died in the presence of all his brethren.

quae respicit Aegyptum introeuntibus Assyrios coram cunctis fratribus suis obiit 25:19 hae quoque sunt generationes 25:19 And these are the generations of Isaac, Abraham’s

son: Abraham begat Isaac:

Isaac filii Abraham Abraham genuit Isaac 25:20 qui cum 25:20 And Isaac was forty years old when he took Re-

bekah to wife, the daughter of Bethuel the Syrian of Padanaram, the sister to Laban the Syrian.

Biblia Sacra Vulgata

quadraginta esset annorum duxit uxorem Rebeccam filiam

25:21 And Isaac intreated the LORD for his wife, because Bathuel Syri de Mesopotamiam sororem Laban 25:21 deshe was barren: and the LORD was intreated of him, and Rebekah his wife conceived.

precatusque est Dominum pro uxore sua eo quod esset ste-

25:22 And the children struggled together within her; and rilis qui exaudiuit eum et dedit conceptum Rebeccae 25:22 she said, If it be so, why am I thus? And she went to enquire of the LORD.

sed conlidebantur in utero eius paruuli quae ait si sic mihi futurum erat quid necesse fuit concipere perrexitque ut con-

25:23 And the LORD said unto her, Two nations are in thy suleret Dominum 25:23 qui respondens ait duae gentes in

womb, and two manner of people shall be separated from thy bowels; and the one people shall be stronger than the

other people; and the elder shall serve the younger. utero tuo sunt et duo populi ex uentre tuo diuidentur popu-25:24 And when her days to be delivered were fulfilled, lusque populum superabit et maior minori seruiet 25:24 iam

behold, there were twins in her womb.

tempus pariendi uenerat et ecce gemini in utero repperti sunt

25:25 And the first came out red, all over like an hairy 25:25 qui primus egressus est rufus erat et totus in morem pel-

garment; and they called his name Esau.

lis hispidus uocatumque est nomen eius Esau protinus alter egrediens plantam fratris tenebat manu et idcirco appellauit

25:26 And after that came his brother out, and his hand eum Iacob 25:26 sexagenarius erat Isaac quando nati sunt

took hold on Esau’s heel; and his name was called Jacob: and Isaac was threescore years old when she bare them.

paruuli 25:27 quibus adultis factus est Esau uir gnarus ue-

25:27 And the boys grew: and Esau was a cunning hunter, a man of the field; and Jacob was a plain man, dwelling in tents.

nandi et homo agricola Iacob autem uir simplex habitabat in

25:28 And Isaac loved Esau, because he did eat of his tabernaculis 25:28 Isaac amabat Esau eo quod de uenationi-

venison: but Rebekah loved Jacob.

25:29 And Jacob sod pottage: and Esau came from the bus illius uesceretur et Rebecca diligebat Iacob 25:29 coxit

field, and he was faint:

autem Iacob pulmentum ad quem cum uenisset Esau de agro

25:30 And Esau said to Jacob, Feed me, I pray thee, with lassus 25:30 ait da mihi de coctione hac rufa quia oppido

that same red pottage; for I am faint: therefore was his name called Edom.

lassus sum quam ob causam uocatum est nomen eius Edom

25:31 And Jacob said, Sell me this day thy birthright. 25:31 cui dixit Iacob uende mihi primogenita tua 25:32 ille

25:32 And Esau said, Behold, I am at the point to die: and what profit shall this birthright do to me?

respondit en morior quid mihi prodeeunt primogenita 25:33

25:33 And Jacob said, Swear to me this day; and he sware unto him: and he sold his birthright unto Jacob.

ait Iacob iura ergo mihi iurauit Esau et uendidit primogenita

25:34 Then Jacob gave Esau bread and pottage of lentiles; 25:34 et sic accepto pane et lentis edulio comedit et bibit et

and he did eat and drink, and rose up, and went his way: thus Esau despised his birthright.

abiit paruipendens quod primogenita uendidisset

26:1 And there was a famine in the land, beside the first 26:1 orta autem fame super terram post eam sterilitatem

famine that was in the days of Abraham. And Isaac went unto Abimelech king of the Philistines unto Gerar.

quae acciderat in diebus Abraham abiit Isaac ad Abimelech

26:2 And the LORD appeared unto him, and said, Go not regem Palestinorum in Gerara 26:2 apparuitque ei Domi-

down into Egypt; dwell in the land which I shall tell thee of:

nus et ait ne descendas in Aegyptum sed quiesce in terra

26:3 Sojourn in this land, and I will be with thee, and will quam dixero tibi 26:3 et peregrinare in ea eroque tecum et bless thee; for unto thee, and unto thy seed, I will give all these countries, and I will perform the oath which I sware

unto Abraham thy father; benedicam tibi tibi enim et semini tuo dabo uniuersas regiones has conplens iuramentum quod spopondi Abraham patri

Liber Genesis

tuo 26:4 et multiplicabo semen tuum sicut stellas caeli da- 26:4 And I will make thy seed to multiply as the stars of

heaven, and will give unto thy seed all these countries; and in thy seed shall all the nations of the earth be blessed;

boque posteris tuis uniuersas regiones has et benedicentur in semine tuo omnes gentes terrae 26:5 eo quod oboedierit 26:5 Because that Abraham obeyed my voice, and kept my

charge, my commandments, my statutes, and my laws.

Abraham uoci meae et custodierit praecepta et mandata mea et caerimonias legesque seruauerit 26:6 mansit itaque Isaac 26:6 And Isaac dwelt in Gerar: in Geraris 26:7 qui cum interrogaretur a uiris loci illius su- 26:7 And the men of the place asked him of his wife; and

he said, She is my sister: for he feared to say, She is my wife; lest, said he, the men of the place should kill me for

per uxore sua respondit soror mea est timuerat enim confi- Rebekah; because she was fair to look upon. teri quod sibi esset sociata coniugio reputans ne forte interficerent eum propter illius pulchritudinem 26:8 cumque per- 26:8 And it came to pass, when he had been there a long

time, that Abimelech king of the Philistines looked out at a window, and saw, and, behold, Isaac was sporting with

transissent dies plurimi et ibi demoraretur prospiciens Abi- Rebekah his wife. melech Palestinorum rex per fenestram uidit eum iocantem cum Rebecca uxore sua 26:9 et accersito ait perspicuum est 26:9 And Abimelech called Isaac, and said, Behold, of a

surety she is thy wife; and how saidst thou, She is my sister? And Isaac said unto him, Because I said, Lest I die

quod uxor tua sit cur mentitus es sororem tuam esse respon- for her. dit timui ne morerer propter eam 26:10 dixitque Abimelech 26:10 And Abimelech said, What is this thou hast done

unto us? one of the people might lightly have lien with thy wife, and thou shouldest have brought guiltiness upon us.

quare inposuisti nobis potuit coire quispiam de populo cum uxore tua et induxeras super nos grande peccatum praecepitque omni populo dicens 26:11 qui tetigerit hominis huius 26:11 And Abimelech charged all his people, saying, He

that toucheth this man or his wife shall surely be put to death.

uxorem morte morietur 26:12 seruit autem Isaac in terra

26:12 Then Isaac sowed in that land, and received in the same year an hundredfold: and the LORD blessed him.

illa et inuenit in ipso anno centuplum benedixitque ei Dominus 26:13 et locupletatus est homo et ibat proficiens atque 26:13 And the man waxed great, and went forward, and

grew until he became very great:

succrescens donec magnus uehementer effectus est 26:14 26:14 For he had possession of flocks, and possession of

herds, and great store of servants: and the Philistines envied him.

habuit quoque possessionem ouium et armentorum et familiae plurimum ob haec inuidentes ei Palestini 26:15 omnes 26:15 For all the wells which his father’s servants had

digged in the days of Abraham his father, the Philistines had stopped them, and filled them with earth.

puteos quos foderant serui patris illius Abraham illo tempore obstruxerunt implentes humo 26:16 in tantum ut ipse 26:16 And Abimelech said unto Isaac, Go from us; for

thou art much mightier than we.

Abimelech diceret ad Isaac recede a nobis quoniam potentior nostri factus es ualde 26:17 et ille discedens ueniret ad 26:17 And Isaac departed thence, and pitched his tent in

the valley of Gerar, and dwelt there.

torrentem Gerarae habitaretque ibi 26:18 rursum fodit alios 26:18 And Isaac digged again the wells of water, which

they had digged in the days of Abraham his father; for the Philistines had stopped them after the death of Abraham: and he called their names after the names by which his father had called them.

puteos quos foderant serui patris sui Abraham et quos illo mortuo olim obstruxerant Philisthim appellauitque eos hisdem nominibus quibus ante pater uocauerat 26:19 foderunt 26:19 And Isaac’s servants digged in the valley, and found

there a well of springing water.

in torrente et reppererunt aquam uiuam 26:20 sed et ibi iur- 26:20 And the herdmen of Gerar did strive with Isaac’s

herdmen, saying, The water is ours: and he called the name of the well Esek; because they strove with him.

gium fuit pastorum Gerarae aduersum pastores Isaac dicen-

Biblia Sacra Vulgata

tium nostra est aqua quam ob rem nomen putei ex eo quod

26:21 And they digged another well, and strove for that acciderat uocauit Calumniam 26:21 foderunt et alium et pro

also: and he called the name of it Sitnah.

26:22 And he removed from thence, and digged another illo quoque rixati sunt appellauitque eum Inimicitias 26:22

well; and for that they strove not: and he called the name of it Rehoboth; and he said, For now the LORD hath made

room for us, and we shall be fruitful in the land. profectus inde fodit alium puteum pro quo non contenderunt

itaque uocauit nomen illius Latitudo dicens nunc dilatauit

26:23 And he went up from thence to Beersheba. nos Dominus et fecit crescere super terram 26:23 ascendit 26:24 And the LORD appeared unto him the same night, autem ex illo loco in Bersabee 26:24 ubi apparuit ei Domi-

and said, I am the God of Abraham thy father: fear not, for I am with thee, and will bless thee, and multiply thy seed

for my servant Abraham’s sake. nus in ipsa nocte dicens ego sum Deus Abraham patris tui

noli metuere quia tecum sum benedicam tibi et multiplicabo

26:25 And he builded an altar there, and called upon the semen tuum propter seruum meum Abraham 26:25 itaque name of the LORD, and pitched his tent there: and there Isaac’s servants digged a well.

aedificauit ibi altare et inuocato nomine Domini extendit tabernaculum praecepitque seruis suis ut foderent puteum

26:26 Then Abimelech went to him from Gerar, and Ahuz- 26:26 ad quem locum cum uenissent de Geraris Abimelech

zath one of his friends, and Phichol the chief captain of his army.

et Ochozath amicus illius et Fichol dux militum 26:27 locu-

26:27 And Isaac said unto them, Wherefore come ye to me, seeing ye hate me, and have sent me away from you?

tus est eis Isaac quid uenistis ad me hominem quem odistis

26:28 And they said, We saw certainly that the LORD was et expulistis a uobis 26:28 qui responderunt uidimus tecum

with thee: and we said, Let there be now an oath betwixt us, even betwixt us and thee, and let us make a covenant

with thee; esse Dominum et idcirco nunc diximus sit iuramentum inter 26:29 That thou wilt do us no hurt, as we have not touched nos et ineamus foedus 26:29 ut non facias nobis quicquam

thee, and as we have done unto thee nothing but good, and have sent thee away in peace: thou art now the blessed of

the LORD. mali sicut et nos nihil tuorum adtigimus nec fecimus quod te

laederet sed cum pace dimisimus auctum benedictione Do-

26:30 And he made them a feast, and they did eat and mini 26:30 fecit ergo eis conuiuium et post cibum et potum

drink.

26:31 And they rose up betimes in the morning, and sware 26:31 surgentes mane iurauerunt sibi mutuo dimisitque eos

one to another: and Isaac sent them away, and they departed from him in peace.

Isaac pacifice in locum suum 26:32 ecce autem uenerunt in

26:32 And it came to pass the same day, that Isaac’s servants came, and told him concerning the well which they had digged, and said unto him, We have found water.

ipso die serui Isaac adnuntiantes ei de puteo quem foderant

26:33 And he called it Shebah: therefore the name of the atque dicentes inuenimus aquam 26:33 unde appellauit eum

city is Beersheba unto this day.

Abundantiam et nomen urbi inpositum est Bersabee usque

26:34 And Esau was forty years old when he took to wife in praesentem diem 26:34 Esau uero quadragenarius duxit

Judith the daughter of Beeri the Hittite, and Bashemath the daughter of Elon the Hittite:

uxores Iudith filiam Beeri Hetthei et Basemath filiam Helon

26:35 Which were a grief of mind unto Isaac and to Re- eiusdem loci 26:35 quae ambae offenderant animum Isaac

bekah.

et Rebeccae

27:1 And it came to pass, that when Isaac was old, and his 27:1 senuit autem Isaac et caligauerunt oculi eius et uidere

eyes were dim, so that he could not see, he called Esau his eldest son, and said unto him, My son: and he said unto

him, Behold, here am I. non poterat uocauitque Esau filium suum maiorem et dixit ei 27:2 And he said, Behold now, I am old, I know not the fili mi qui respondit adsum 27:2 cui pater uides inquit quod

day of my death:

Liber Genesis

senuerim et ignorem diem mortis meae 27:3 sume arma tua 27:3 Now therefore take, I pray thee, thy weapons, thy

quiver and thy bow, and go out to the field, and take me some venison;

faretram et arcum et egredere foras cumque uenatu aliquid adprehenderis 27:4 fac mihi inde pulmentum sicut uelle me 27:4 And make me savoury meat, such as I love, and bring

it to me, that I may eat; that my soul may bless thee before I die.

nosti et adfer ut comedam et benedicat tibi anima mea antequam moriar 27:5 quod cum audisset Rebecca et ille abis- 27:5 And Rebekah heard when Isaac spake to Esau his

son. And Esau went to the field to hunt for venison, and to bring it.

set in agrum ut iussionem patris expleret 27:6 dixit filio suo

27:6 And Rebekah spake unto Jacob her son, saying, Behold, I heard thy father speak unto Esau thy brother, saying,

Iacob audiui patrem tuum loquentem cum Esau fratre tuo et dicentem ei 27:7 adfer mihi uenationem tuam et fac cibos 27:7 Bring me venison, and make me savoury meat, that I

may eat, and bless thee before the LORD before my death.

ut comedam et benedicam tibi coram Domino antequam moriar 27:8 nunc ergo fili mi adquiesce consiliis meis 27:9 et 27:8 Now therefore, my son, obey my voice according to

that which I command thee.

27:9 Go now to the flock, and fetch me from thence two good kids of the goats; and I will make them savoury meat for thy father, such as he loveth:

pergens ad gregem adfer mihi duos hedos optimos ut faciam ex eis escas patri tuo quibus libenter uescitur 27:10 quas 27:10 And thou shalt bring it to thy father, that he may eat,

and that he may bless thee before his death.

cum intuleris et comederit benedicat tibi priusquam moriatur 27:11 cui ille respondit nosti quod Esau frater meus homo pi- 27:11 And Jacob said to Rebekah his mother, Behold,

Esau my brother is a hairy man, and I am a smooth man:

losus sit et ego lenis 27:12 si adtractauerit me pater meus 27:12 My father peradventure will feel me, and I shall

seem to him as a deceiver; and I shall bring a curse upon me, and not a blessing.

et senserit timeo ne putet sibi uoluisse inludere et inducat super me maledictionem pro benedictione 27:13 ad quem 27:13 And his mother said unto him, Upon me be thy

curse, my son: only obey my voice, and go fetch me them.

mater in me sit ait ista maledictio fili mi tantum audi uocem meam et perge adferque quae dixi 27:14 abiit et adtulit 27:14 And he went, and fetched, and brought them to his

mother: and his mother made savoury meat, such as his father loved.

deditque matri parauit illa cibos sicut nouerat uelle patrem illius 27:15 et uestibus Esau ualde bonis quas apud se habe- 27:15 And Rebekah took goodly raiment of her eldest son

Esau, which were with her in the house, and put them upon Jacob her younger son:

bat domi induit eum 27:16 pelliculasque hedorum circum-

27:16 And she put the skins of the kids of the goats upon his hands, and upon the smooth of his neck:

dedit manibus et colli nuda protexit 27:17 dedit pulmentum _{27:17 And she gave the savoury meat and the bread, which}

she had prepared, into the hand of her son Jacob.

et panes quos coxerat tradidit 27:18 quibus inlatis dixit pa- 27:18 And he came unto his father, and said, My father:

and he said, Here am I; who art thou, my son?

ter mi et ille respondit audio quis tu es fili mi 27:19 dixit- 27:19 And Jacob said unto his father, I am Esau thy first

born; I have done according as thou badest me: arise, I pray thee, sit and eat of my venison, that thy soul may

que Iacob ego sum Esau primogenitus tuus feci sicut prae- bless me. cepisti mihi surge sede et comede de uenatione mea ut be-

nedicat mihi anima tua 27:20 rursum Isaac ad filium suum 27:20 And Isaac said unto his son, How is it that thou hast found it so quickly, my son? And he said, Because the LORD thy God brought it to me.

quomodo inquit tam cito inuenire potuisti fili mi qui respondit uoluntatis Dei fuit ut cito mihi occurreret quod uolebam 27:21 dixitque Isaac accede huc ut tangam te fili mi et pro- 27:21 And Isaac said unto Jacob, Come near, I pray thee,

that I may feel thee, my son, whether thou be my very son Esau or not.

bem utrum tu sis filius meus Esau an non 27:22 accessit ille

27:22 And Jacob went near unto Isaac his father; and he felt him, and said, The voice is Jacob’s voice, but the hands are the hands of Esau.

ad patrem et palpato eo dixit Isaac uox quidem uox Iacob

Biblia Sacra Vulgata

27:23 And he discerned him not, because his hands were est sed manus manus sunt Esau 27:23 et non cognouit eum hairy, as his brother Esau’s hands: so he blessed him.

quia pilosae manus similitudinem maioris expresserant be-

27:24 And he said, Art thou my very son Esau? And he nedicens ergo illi 27:24 ait tu es filius meus Esau respondit

said, I am.

27:25 And he said, Bring it near to me, and I will eat of ego sum 27:25 at ille offer inquit mihi cibos de uenatione

my son’s venison, that my soul may bless thee. And he brought it near to him, and he did eat: and he brought him

wine and he drank. tua fili mi ut benedicat tibi anima mea quos cum oblatos co-27:26 And his father Isaac said unto him, Come near now, medisset obtulit ei etiam uinum quo hausto 27:26 dixit ad

and kiss me, my son.

27:27 And he came near, and kissed him: and he smelled eum accede ad me et da mihi osculum fili mi 27:27 accessit

the smell of his raiment, and blessed him, and said, See, the smell of my son is as the smell of a field which the

LORD hath blessed: et osculatus est eum statimque ut sensit uestimentorum illius

flagrantiam benedicens ait ecce odor filii mei sicut odor agri

27:28 Therefore God give thee of the dew of heaven, and cui benedixit Dominus 27:28 det tibi Deus de rore caeli et

the fatness of the earth, and plenty of corn and wine:

27:29 Let people serve thee, and nations bow down to de pinguedine terrae abundantiam frumenti et uini 27:29 et

thee: be lord over thy brethren, and let thy mother’s sons bow down to thee: cursed be every one that curseth thee,

and blessed be he that blesseth thee. seruiant tibi populi et adorent te tribus esto dominus fratrum tuorum et incuruentur ante te filii matris tuae qui maledixerit tibi sit maledictus et qui benedixerit benedictionibus reple-

27:30 And it came to pass, as soon as Isaac had made an atur 27:30 uix Isaac sermonem impleuerat et egresso Iacob

end of blessing Jacob, and Jacob was yet scarce gone out from the presence of Isaac his father, that Esau his brother

came in from his hunting. foras uenit Esau 27:31 coctosque de uenatione cibos intulit

27:31 And he also had made savoury meat, and brought it unto his father, and said unto his father, Let my father arise, and eat of his son’s venison, that thy soul may bless me.

patri dicens surge pater mi et comede de uenatione filii tui

27:32 And Isaac his father said unto him, Who art thou? ut benedicat mihi anima tua 27:32 dixitque illi Isaac quis

And he said, I am thy son, thy firstborn Esau.

enim es tu qui respondit ego sum primogenitus filius tuus

27:33 And Isaac trembled very exceedingly, and said, Esau 27:33 expauit Isaac stupore uehementi et ultra quam

Who? where is he that hath taken venison, and brought it me, and I have eaten of all before thou camest, and have

blessed him? yea, and he shall be blessed. credi potest admirans ait quis igitur ille est qui dudum ca-

ptam uenationem adtulit mihi et comedi ex omnibus prius-

27:34 And when Esau heard the words of his father, he quam tu uenires benedixique ei et erit benedictus 27:34 au-

cried with a great and exceeding bitter cry, and said unto his father, Bless me, even me also, O my father.

ditis Esau sermonibus patris inrugiit clamore magno et con-

27:35 And he said, Thy brother came with subtilty, and sternatus ait benedic etiam mihi pater mi 27:35 qui ait uenit

hath taken away thy blessing.

germanus tuus fraudulenter et accepit benedictionem tuam

27:36 And he said, Is not he rightly named Jacob? for 27:36 at ille subiunxit iuste uocatum est nomen eius Iacob

he hath supplanted me these two times: he took away my birthright; and, behold, now he hath taken away my blessing. And he said, Hast thou not reserved a blessing for me?

subplantauit enim me en altera uice primogenita mea ante tulit et nunc secundo subripuit benedictionem meam rursumque ad patrem numquid non reseruasti ait et mihi benedic-

27:37 And Isaac answered and said unto Esau, Behold, I tionem 27:37 respondit Isaac dominum tuum illum consti-

have made him thy lord, and all his brethren have I given to him for servants; and with corn and wine have I sustained

him: and what shall I do now unto thee, my son? tui et omnes fratres eius seruituti illius subiugaui frumento et uino stabiliui eum tibi post haec fili mi ultra quid faciam

Liber Genesis

27:38 cui Esau num unam inquit tantum benedictionem ha- 27:38 And Esau said unto his father, Hast thou but one

blessing, my father? bless me, even me also, O my father. And Esau lifted up his voice, and wept.

bes pater mihi quoque obsecro ut benedicas cumque heiulatu magno fleret 27:39 motus Isaac dixit ad eum in pin- 27:39 And Isaac his father answered and said unto him,

Behold, thy dwelling shall be the fatness of the earth, and of the dew of heaven from above;

guedine terrae et in rore caeli desuper 27:40 erit benedictio

27:40 And by thy sword shalt thou live, and shalt serve thy brother; and it shall come to pass when thou shalt have the dominion, that thou shalt break his yoke from off thy neck.

tua uiues gladio et fratri tuo seruies tempusque ueniet cum excutias et soluas iugum eius de ceruicibus tuis 27:41 ode- 27:41 And Esau hated Jacob because of the blessing

wherewith his father blessed him: and Esau said in his heart, The days of mourning for my father are at hand;

rat ergo semper Esau Iacob pro benedictione qua benedixerat then will I slay my brother Jacob. ei pater dixitque in corde suo ueniant dies luctus patris mei ut occidam Iacob fratrem meum 27:42 nuntiata sunt haec 27:42 And these words of Esau her elder son were told to

Rebekah: and she sent and called Jacob her younger son, and said unto him, Behold, thy brother Esau, as touching

Rebeccae quae mittens et uocans Iacob filium suum dixit ad thee, doth comfort himself, purposing to kill thee. eum ecce Esau frater tuus minatur ut occidat te 27:43 nunc 27:43 Now therefore, my son, obey my voice; arise, flee

thou to Laban my brother to Haran;

ergo fili audi uocem meam et consurgens fuge ad Laban fratrem meum in Haran 27:44 habitabisque cum eo dies paucos 27:44 And tarry with him a few days, until thy brother’s

fury turn away;

donec requiescat furor fratris tui 27:45 et cesset indignatio 27:45 Until thy brother’s anger turn away from thee, and

he forget that which thou hast done to him: then I will send, and fetch thee from thence: why should I be de-

eius obliuiscaturque eorum quae fecisti in eum postea mit- prived also of you both in one day? tam et adducam te inde huc cur utroque orbabor filio in una die 27:46 dixit quoque Rebecca ad Isaac taedet me uitae 27:46 And Rebekah said to Isaac, I am weary of my life

because of the daughters of Heth: if Jacob take a wife of the daughters of Heth, such as these which are of the

meae propter filias Heth si acceperit Iacob uxorem de stirpe daughters of the land, what good shall my life do me? huius terrae nolo uiuere

28:1 uocauit itaque Isaac Iacob et benedixit praecepitque 28:1 And Isaac called Jacob, and blessed him, and charged

him, and said unto him, Thou shalt not take a wife of the daughters of Canaan.

ei dicens noli accipere coniugem de genere Chanaan 28:2

28:2 Arise, go to Padanaram, to the house of Bethuel thy mother’s father; and take thee a wife from thence of the daughers of Laban thy mother’s brother.

sed uade et proficiscere in Mesopotamiam Syriae ad domum Bathuel patrem matris tuae et accipe tibi inde uxo-

rem de filiabus Laban auunculi tui 28:3 Deus autem om- 28:3 And God Almighty bless thee, and make thee fruitful, and multiply thee, that thou mayest be a multitude of people;

nipotens benedicat tibi et crescere te faciat atque multiplicet ut sis in turbas populorum 28:4 et det tibi benedic- 28:4 And give thee the blessing of Abraham, to thee, and

to thy seed with thee; that thou mayest inherit the land wherein thou art a stranger, which God gave unto Abra-

tiones Abraham et semini tuo post te ut possideas terram ham. peregrinationis tuae quam pollicitus est auo tuo 28:5 cum- 28:5 And Isaac sent away Jacob: and he went to Pada-

naram unto Laban, son of Bethuel the Syrian, the brother of Rebekah, Jacob’s and Esau’s mother.

que dimisisset eum Isaac profectus uenit in Mesopotamiam Syriae ad Laban filium Bathuel Syri fratrem Rebeccae matris suae 28:6 uidens autem Esau quod benedixisset pater 28:6 When Esau saw that Isaac had blessed Jacob, and sent

him away to Padanaram, to take him a wife from thence; and that as he blessed him he gave him a charge, saying,

suus Iacob et misisset eum in Mesopotamiam Syriae ut in- Thou shalt not take a wife of the daughers of Canaan; de uxorem duceret et quod post benedictionem praecepisset

Biblia Sacra Vulgata

28:7 And that Jacob obeyed his father and his mother, and ei dicens non accipies coniugem de filiabus Chanaan 28:7 was gone to Padanaram;

28:8 And Esau seeing that the daughters of Canaan quodque oboediens Iacob parentibus isset in Syriam 28:8

pleased not Isaac his father;

probans quoque quod non libenter aspiceret filias Chanaan

28:9 Then went Esau unto Ishmael, and took unto the pater suus 28:9 iuit ad Ismahelem et duxit uxorem absque

wives which he had Mahalath the daughter of Ishmael Abraham’s son, the sister of Nebajoth, to be his wife.

his quas prius habebat Maeleth filiam Ismahel filii Abraham

28:10 And Jacob went out from Beersheba, and went to- sororem Nabaioth 28:10 igitur egressus Iacob de Bersabee

ward Haran.

28:11 And he lighted upon a certain place, and tarried pergebat Haran 28:11 cumque uenisset ad quendam locum

there all night, because the sun was set; and he took of the stones of that place, and put them for his pillows, and

lay down in that place to sleep. et uellet in eo requiescere post solis occubitum tulit de lapi-

dibus qui iacebant et subponens capiti suo dormiuit in eodem

28:12 And he dreamed, and behold a ladder set up on the loco 28:12 uiditque in somnis scalam stantem super terram

earth, and the top of it reached to heaven: and behold the angels of God ascending and descending on it.

et cacumen illius tangens caelum angelos quoque Dei ascen-

28:13 And, behold, the LORD stood above it, and said, I dentes et descendentes per eam 28:13 et Dominum innixum

am the LORD God of Abraham thy father, and the God of Isaac: the land whereon thou liest, to thee will I give it,

and to thy seed; scalae dicentem sibi ego sum Dominus Deus Abraham patris

tui et Deus Isaac terram in qua dormis tibi dabo et semini tuo

28:14 And thy seed shall be as the dust of the earth, and 28:14 eritque germen tuum quasi puluis terrae dilataberis ad

thou shalt spread abroad to the west, and to the east, and to the north, and to the south: and in thee and in thy seed

shall all the families of the earth be blessed. occidentem et orientem septentrionem et meridiem et bene-28:15 And, behold, I am with thee, and will keep thee in dicentur in te et in semine tuo cunctae tribus terrae 28:15 et

all places whither thou goest, and will bring thee again into this land; for I will not leave thee, until I have done

that which I have spoken to thee of. ero custos tuus quocumque perrexeris et reducam te in ter-

ram hanc nec dimittam nisi conpleuero uniuersa quae dixi

28:16 And Jacob awaked out of his sleep, and he said, 28:16 cumque euigilasset Iacob de somno ait uere Dominus

Surely the LORD is in this place; and I knew it not.

28:17 And he was afraid, and said, How dreadful is this est in loco isto et ego nesciebam 28:17 pauensque quam ter-

place! this is none other but the house of God, and this is the gate of heaven.

ribilis inquit est locus iste non est hic aliud nisi domus Dei

28:18 And Jacob rose up early in the morning, and took et porta caeli 28:18 surgens ergo mane tulit lapidem quem

the stone that he had put for his pillows, and set it up for a pillar, and poured oil upon the top of it.

subposuerat capiti suo et erexit in titulum fundens oleum de-

28:19 And he called the name of that place Bethel: but the super 28:19 appellauitque nomen urbis Bethel quae prius

name of that city was called Luz at the first.

28:20 And Jacob vowed a vow, saying, If God will be with Luza uocabatur 28:20 uouit etiam uotum dicens si fuerit

me, and will keep me in this way that I go, and will give me bread to eat, and raiment to put on,

Deus mecum et custodierit me in uia per quam ambulo et dederit mihi panem ad uescendum et uestem ad induendum

28:21 So that I come again to my father’s house in peace; 28:21 reuersusque fuero prospere ad domum patris mei erit

then shall the LORD be my God: 28:22 And this stone, which I have set for a pillar, shall mihi Dominus in Deum 28:22 et lapis iste quem erexi in be God’s house: and of all that thou shalt give me I will surely give the tenth unto thee.

titulum uocabitur Domus Dei cunctorumque quae dederis mihi decimas offeram tibi

29:1 Then Jacob went on his journey, and came into the 29:1 profectus ergo Iacob uenit ad terram orientalem 29:2

land of the people of the east. 29:2 And he looked, and behold a well in the field, and, lo, there were three flocks of sheep lying by it; for out of that well they watered the flocks: and a great stone was upon the well’s mouth.

XLVI

Liber Genesis

et uidit puteum in agro tresque greges ouium accubantes iuxta eum nam ex illo adaquabantur pecora et os eius grandi lapide claudebatur 29:3 morisque erat ut cunctis ouibus 29:3 And thither were all the flocks gathered: and they

rolled the stone from the well’s mouth, and watered the sheep, and put the stone again upon the well’s mouth in

congregatis deuoluerent lapidem et refectis gregibus rursum his place. super os putei ponerent 29:4 dixitque ad pastores fratres 29:4 And Jacob said unto them, My brethren, whence be

ye? And they said, Of Haran are we.

unde estis qui responderunt de Haran 29:5 quos interrogans 29:5 And he said unto them, Know ye Laban the son of

Nahor? And they said, We know him.

numquid ait nostis Laban filium Nahor dixerunt nouimus 29:6 sanusne est inquit ualet inquiunt et ecce Rahel filia eius 29:6 And he said unto them, Is he well? And they said, He

is well: and, behold, Rachel his daughter cometh with the sheep.

uenit cum grege suo 29:7 dixitque Iacob adhuc multum diei

29:7 And he said, Lo, it is yet high day, neither is it time that the cattle should be gathered together: water ye the sheep, and go and feed them.

superest nec est tempus ut reducantur ad caulas greges date

ante potum ouibus et sic ad pastum eas reducite 29:8 qui 29:8 And they said, We cannot, until all the flocks be gathered together, and till they roll the stone from the well’s mouth; then we water the sheep.

responderunt non possumus donec omnia pecora congregentur et amoueamus lapidem de ore putei ut adaquemus greges 29:9 adhuc loquebantur et ecce Rahel ueniebat cum ouibus 29:9 And while he yet spake with them, Rachel came with

her father’s sheep; for she kept them.

patris sui nam gregem ipsa pascebat 29:10 quam cum uidis- 29:10 And it came to pass, when Jacob saw Rachel the

daughter of Laban his mother’s brother, and the sheep of Laban his mother’s brother, that Jacob went near, and rolled the stone from the well’s mouth, and watered the flock of Laban his mother’s brother.

set Iacob et sciret consobrinam suam ouesque Laban auunculi sui amouit lapidem quo puteus claudebatur 29:11 et ad- 29:11 And Jacob kissed Rachel, and lifted up his voice,

and wept.

aquato grege osculatus est eam eleuataque uoce fleuit 29:12 29:12 And Jacob told Rachel that he was her father’s

brother, and that he was Rebekah’s son: and she ran and told her father.

et indicauit ei quod frater esset patris eius et filius Rebeccae at illa festinans nuntiauit patri suo 29:13 qui cum audisset 29:13 And it came to pass, when Laban heard the tidings

of Jacob his sister’s son, that he ran to meet him, and embraced him, and kissed him, and brought him to his house.

uenisse Iacob filium sororis suae cucurrit obuiam conplexus- And he told Laban all these things. que eum et in oscula ruens duxit in domum suam auditis autem causis itineris 29:14 respondit os meum es et caro mea 29:14 And Laban said to him, Surely thou art my bone and

my flesh. And he abode with him the space of a month.

et postquam expleti sunt dies mensis unius 29:15 dixit ei 29:15 And Laban said unto Jacob, Because thou art my

brother, shouldest thou therefore serve me for nought? tell me, what shall thy wages be?

num quia frater meus es gratis seruies mihi dic quid mercedis accipias 29:16 habebat uero filias duas nomen maio- 29:16 And Laban had two daughters: the name of the elder

was Leah, and the name of the younger was Rachel.

ris Lia minor appellabatur Rahel 29:17 sed Lia lippis erat 29:17 Leah was tender eyed; but Rachel was beautiful and

well favoured.

oculis Rahel decora facie et uenusto aspectu 29:18 quam 29:18 And Jacob loved Rachel; and said, I will serve thee

seven years for Rachel thy younger daughter.

diligens Iacob ait seruiam tibi pro Rahel filia tua minore se-

ptem annis 29:19 respondit Laban melius est ut tibi eam 29:19 And Laban said, It is better that I give her to thee,

than that I should give her to another man: abide with me.

dem quam uiro alteri mane apud me 29:20 seruiuit igitur Ia- 29:20 And Jacob served seven years for Rachel; and they

seemed unto him but a few days, for the love he had to her.

cob pro Rahel septem annis et uidebantur illi pauci dies prae amoris magnitudine 29:21 dixitque ad Laban da mihi uxo- 29:21 And Jacob said unto Laban, Give me my wife, for

my days are fulfilled, that I may go in unto her.

Biblia Sacra Vulgata

rem meam quia iam tempus expletum est ut ingrediar ad eam

29:22 And Laban gathered together all the men of the 29:22 qui uocatis multis amicorum turbis ad conuiuium fe-

place, and made a feast.

29:23 And it came to pass in the evening, that he took Leah cit nuptias 29:23 et uespere filiam suam Liam introduxit ad

his daughter, and brought her to him; and he went in unto her.

eum 29:24 dans ancillam filiae Zelpham nomine ad quam

29:24 And Laban gave unto his daughter Leah Zilpah his maid for an handmaid.

cum ex more Iacob fuisset ingressus facto mane uidit Liam

29:25 And it came to pass, that in the morning, behold, 29:25 et dixit ad socerum quid est quod facere uoluisti nonne

it was Leah: and he said to Laban, What is this thou hast done unto me? did not I serve with thee for Rachel?

wherefore then hast thou beguiled me? pro Rahel seruiui tibi quare inposuisti mihi 29:26 respon-

29:26 And Laban said, It must not be so done in our country, to give the younger before the firstborn.

dit Laban non est in loco nostro consuetudinis ut minores

29:27 Fulfil her week, and we will give thee this also for ante tradamus ad nuptias 29:27 imple ebdomadem dierum the service which thou shalt serve with me yet seven other years.

huius copulae et hanc quoque dabo tibi pro opere quo ser-

29:28 And Jacob did so, and fulfilled her week: and he uiturus es mihi septem annis aliis 29:28 adquieuit placito

gave him Rachel his daughter to wife also.

29:29 And Laban gave to Rachel his daughter Bilhah his et ebdomade transacta Rahel duxit uxorem 29:29 cui pater

handmaid to be her maid.

29:30 And he went in also unto Rachel, and he loved also seruam Balam dederat 29:30 tandemque potitus optatis nu-

Rachel more than Leah, and served with him yet seven other years.

ptiis amorem sequentis priori praetulit seruiens apud eum

29:31 And when the LORD saw that Leah was hated, he septem annis aliis 29:31 uidens autem Dominus quod de-

opened her womb: but Rachel was barren.

spiceret Liam aperuit uuluam eius sorore sterili permanente

29:32 And Leah conceived, and bare a son, and she called 29:32 quae conceptum genuit filium uocauitque nomen eius

his name Reuben: for she said, Surely the LORD hath looked upon my affliction; now therefore my husband will

love me. Ruben dicens uidit Dominus humilitatem meam nunc ama-29:33 And she conceived again, and bare a son; and said, bit me uir meus 29:33 rursumque concepit et peperit filium

Because the LORD hath heard I was hated, he hath therefore given me this son also: and she called his name

Simeon. et ait quoniam audiuit Dominus haberi me contemptui de-29:34 And she conceived again, and bare a son; and said, dit etiam istum mihi uocauitque nomen illius Symeon 29:34

Now this time will my husband be joined unto me, because I have born him three sons: therefore was his name called

Levi. concepit tertio et genuit alium dixitque nunc quoque copu-

labitur mihi maritus meus eo quod pepererim illi tres filios

29:35 And she conceived again, and bare a son: and she et idcirco appellauit nomen eius Leui 29:35 quarto conce-

said, Now will I praise the LORD: therefore she called his name Judah; and left bearing.

pit et peperit filium et ait modo confitebor Domino et ob hoc uocauit eum Iudam cessauitque parere

30:1 And when Rachel saw that she bare Jacob no chil- 30:1 cernens autem Rahel quod infecunda esset inuidit so-

dren, Rachel envied her sister; and said unto Jacob, Give me children, or else I die.

rori et ait marito suo da mihi liberos alioquin moriar 30:2

30:2 And Jacob’s anger was kindled against Rachel: and he said, Am I in God’s stead, who hath withheld from thee the fruit of the womb?

cui iratus respondit Iacob num pro Deo ego sum qui priuauit

30:3 And she said, Behold my maid Bilhah, go in unto te fructu uentris tui 30:3 at illa habeo inquit famulam Balam

her; and she shall bear upon my knees, that I may also have children by her.

ingredere ad eam ut pariat super genua mea et habeam ex

30:4 And she gave him Bilhah her handmaid to wife: and ea filios 30:4 deditque illi Balam in coniugium quae 30:5

Jacob went in unto her. 30:5 And Bilhah conceived, and bare Jacob a son.

ingresso ad se uiro concepit et peperit filium 30:6 dixit-

30:6 And Rachel said, God hath judged me, and hath also heard my voice, and hath given me a son: therefore called she his name Dan.

Liber Genesis

que Rahel iudicauit mihi Dominus et exaudiuit uocem meam dans mihi filium et idcirco appellauit nomen illius Dan 30:7 30:7 And Bilhah Rachel’s maid conceived again, and bare

Jacob a second son.

rursumque Bala concipiens peperit alterum 30:8 pro quo ait 30:8 And Rachel said, With great wrestlings have I wres-

tled with my sister, and I have prevailed: and she called his name Naphtali.

Rahel conparauit me Deus cum sorore mea et inualui uocauitque eum Nepthalim 30:9 sentiens Lia quod parere desis- 30:9 When Leah saw that she had left bearing, she took

Zilpah her maid, and gave her Jacob to wife.

set Zelpham ancillam suam marito tradidit 30:10 qua post 30:10 And Zilpah Leah’s maid bare Jacob a son. conceptum edente filium 30:11 dixit feliciter et idcirco uo- 30:11 And Leah said, A troop cometh: and she called his

name Gad.

cauit nomen eius Gad 30:12 peperit quoque Zelpha alterum 30:12 And Zilpah Leah’s maid bare Jacob a second son. 30:13 dixitque Lia hoc pro beatitudine mea beatam quippe me 30:13 And Leah said, Happy am I, for the daughters will

call me blessed: and she called his name Asher.

dicent mulieres propterea appellauit eum Aser 30:14 egres- 30:14 And Reuben went in the days of wheat harvest, and

found mandrakes in the field, and brought them unto his mother Leah. Then Rachel said to Leah, Give me, I pray

sus autem Ruben tempore messis triticeae in agro repperit thee, of thy son’s mandrakes. mandragoras quos matri Liae detulit dixitque Rahel da mihi partem de mandragoris filii tui 30:15 illa respondit parumne 30:15 And she said unto her, Is it a small matter that thou

hast taken my husband? and wouldest thou take away my son’s mandrakes also? And Rachel said, Therefore

tibi uidetur quod praeripueris maritum mihi nisi etiam man- he shall lie with thee to night for thy son’s mandrakes. dragoras filii mei tuleris ait Rahel dormiat tecum hac nocte pro mandragoris filii tui 30:16 redeentique ad uesperam de 30:16 And Jacob came out of the field in the evening, and

Leah went out to meet him, and said, Thou must come in unto me; for surely I have hired thee with my son’s

agro Iacob egressa est in occursum Lia et ad me inquit intra- mandrakes. And he lay with her that night. bis quia mercede conduxi te pro mandragoris filii mei dormiuit cum ea nocte illa 30:17 et exaudiuit Deus preces eius 30:17 And God hearkened unto Leah, and she conceived,

and bare Jacob the fifth son.

concepitque et peperit filium quintum 30:18 et ait dedit De- 30:18 And Leah said, God hath given me my hire, because

I have given my maiden to my husband: and she called his name Issachar.

us mercedem mihi quia dedi ancillam meam uiro meo appellauitque nomen illius Isachar 30:19 rursum Lia concipiens 30:19 And Leah conceived again, and bare Jacob the sixth

son.

peperit sextum filium 30:20 et ait ditauit me Deus dote bona 30:20 And Leah said, God hath endued me with a good

dowry; now will my husband dwell with me, because I have born him six sons: and she called his name Zebulun.

etiam hac uice mecum erit maritus meus eo quod genuerim ei sex filios et idcirco appellauit nomen eius Zabulon 30:21 30:21 And afterwards she bare a daughter, and called her

name Dinah.

post quem peperit filiam nomine Dinam 30:22 recordatus 30:22 And God remembered Rachel, and God hearkened

to her, and opened her womb.

quoque Dominus Rahelis exaudiuit eam et aperuit uuluam illius 30:23 quae concepit et peperit filium dicens abstulit 30:23 And she conceived, and bare a son; and said, God

hath taken away my reproach:

Deus obprobrium meum 30:24 et uocauit nomen illius Io- 30:24 And she called his name Joseph; and said, The

LORD shall add to me another son.

seph dicens addat mihi Dominus filium alterum 30:25 nato 30:25 And it came to pass, when Rachel had born Joseph,

that Jacob said unto Laban, Send me away, that I may go unto mine own place, and to my country.

autem Ioseph dixit Iacob socero suo dimitte me ut reuertar in patriam et ad terram meam 30:26 da mihi uxores et 30:26 Give me my wives and my children, for whom I

have served thee, and let me go: for thou knowest my service which I have done thee.

liberos meos pro quibus seruiui tibi ut abeam tu nosti ser-

Biblia Sacra Vulgata

30:27 And Laban said unto him, I pray thee, if I have uitutem qua seruiui tibi 30:27 ait ei Laban inueniam gra-

found favour in thine eyes, tarry: for I have learned by experience that the LORD hath blessed me for thy sake.

tiam in conspectu tuo experimento didici quod benedixerit

30:28 And he said, Appoint me thy wages, and I will give mihi Deus propter te 30:28 constitue mercedem tuam quam

it.

30:29 And he said unto him, Thou knowest how I have dem tibi 30:29 at ille respondit tu nosti quomodo seruierim

served thee, and how thy cattle was with me.

30:30 For it was little which thou hadst before I came, tibi et quanta in manibus meis fuerit possessio tua 30:30

and it is now increased unto a multitude; and the LORD hath blessed thee since my coming: and now when shall I

provide for mine own house also? modicum habuisti antequam uenirem et nunc diues effec-

tus es benedixitque tibi Dominus ad introitum meum iustum

30:31 And he said, What shall I give thee? And Jacob est igitur ut aliquando prouideam etiam domui meae 30:31

said, Thou shalt not give me any thing: if thou wilt do this thing for me, I will again feed and keep thy flock.

dixitque Laban quid dabo tibi at ille ait nihil uolo sed si feceris quod postulo iterum pascam et custodiam pecora tua

30:32 I will pass through all thy flock to day, removing 30:32 gyra omnes greges tuos et separa cunctas oues uarias

from thence all the speckled and spotted cattle, and all the brown cattle among the sheep, and the spotted and speck-

led among the goats: and of such shall be my hire. et sparso uellere et quodcumque furuum et maculosum ua-

riumque fuerit tam in ouibus quam in capris erit merces mea

30:33 So shall my righteousness answer for me in time 30:33 respondebitque mihi cras iustitia mea quando placiti

to come, when it shall come for my hire before thy face: every one that is not speckled and spotted among the goats, and brown among the sheep, that shall be counted stolen with me.

tempus aduenerit coram te et omnia quae non fuerint uaria et maculosa et furua tam in ouibus quam in capris furti me

30:34 And Laban said, Behold, I would it might be ac- arguent 30:34 dixit Laban gratum habeo quod petis 30:35

cording to thy word. 30:35 And he removed that day the he goats that were ringstraked and spotted, and all the she goats that were speckled and spotted, and every one that had some white in it, and all the brown among the sheep, and gave them into the hand of his sons.

et separauit in die illo capras et oues hircos et arietes uarios atque maculosos cunctum autem gregem unicolorem id est

30:36 And he set three days’ journey betwixt himself and albi et nigri uelleris tradidit in manu filiorum suorum 30:36

Jacob: and Jacob fed the rest of Laban’s flocks.

et posuit spatium itineris inter se et generum dierum trium

30:37 And Jacob took him rods of green poplar, and of the qui pascebat reliquos greges eius 30:37 tollens ergo Iacob

hazel and chesnut tree; and pilled white strakes in them, and made the white appear which was in the rods.

uirgas populeas uirides et amigdalinas et ex platanis ex parte decorticauit eas detractisque corticibus in his quae spoliata fuerant candor apparuit illa uero quae integra erant uiridia permanserunt atque in hunc modum color effectus est ua-

30:38 And he set the rods which he had pilled before rius 30:38 posuitque eas in canalibus ubi effundebatur aqua

the flocks in the gutters in the watering troughs when the flocks came to drink, that they should conceive when they

came to drink. ut cum uenissent greges ad bibendum ante oculos haberent 30:39 And the flocks conceived before the rods, and uirgas et in aspectu earum conciperent 30:39 factumque est

brought forth cattle ringstraked, speckled, and spotted.

ut in ipso calore coitus oues intuerentur uirgas et parerent

30:40 And Jacob did separate the lambs, and set the faces maculosa et uaria et diuerso colore respersa 30:40 diuisit-

of the flocks toward the ringstraked, and all the brown in the flock of Laban; and he put his own flocks by them-

selves, and put them not unto Laban’s cattle. que gregem Iacob et posuit uirgas ante oculos arietum erant autem alba quaeque et nigra Laban cetera uero Iacob separa-

Liber Genesis

tis inter se gregibus 30:41 igitur quando primo tempore as- 30:41 And it came to pass, whensoever the stronger cattle

did conceive, that Jacob laid the rods before the eyes of the cattle in the gutters, that they might conceive among

cendebantur oues ponebat Iacob uirgas in canalibus aquarum the rods. ante oculos arietum et ouium ut in earum contemplatione conciperent 30:42 quando uero serotina admissura erat et 30:42 But when the cattle were feeble, he put them not in:

so the feebler were Laban’s, and the stronger Jacob’s.

conceptus extremus non ponebat eas factaque sunt ea quae erant serotina Laban et quae primi temporis Iacob 30:43 di- 30:43 And the man increased exceedingly, and had much

cattle, and maidservants, and menservants, and camels, and asses.

tatusque est homo ultra modum et habuit greges multos ancillas et seruos camelos et asinos

31:1 postquam autem audiuit uerba filiorum Laban dicen- 31:1 And he heard the words of Laban’s sons, saying, Ja-

cob hath taken away all that was our father’s; and of that which was our father’s hath he gotten all this glory.

tium tulit Iacob omnia quae fuerunt patris nostri et de illius facultate ditatus factus est inclitus 31:2 animaduertit quo- 31:2 And Jacob beheld the countenance of Laban, and,

behold, it was not toward him as before.

que faciem Laban quod non esset erga se sicut heri et nudius tertius 31:3 maxime dicente sibi Domino reuertere in 31:3 And the LORD said unto Jacob, Return unto the land

of thy fathers, and to thy kindred; and I will be with thee.

terram patrum tuorum et ad generationem tuam eroque tecum 31:4 misit et uocauit Rahel et Liam in agrum ubi pa- 31:4 And Jacob sent and called Rachel and Leah to the

field unto his flock,

scebat greges 31:5 dixitque eis uideo faciem patris uestri 31:5 And said unto them, I see your father’s countenance,

that it is not toward me as before; but the God of my father hath been with me.

quod non sit erga me sicut heri et nudius tertius Deus autem patris mei fuit mecum 31:6 et ipsae nostis quod totis uiribus 31:6 And ye know that with all my power I have served

your father.

meis seruierim patri uestro 31:7 sed pater uester circumue- 31:7 And your father hath deceived me, and changed my

wages ten times; but God suffered him not to hurt me.

nit me et mutauit mercedem meam decem uicibus et tamen non dimisit eum Deus ut noceret mihi 31:8 si quando dixit 31:8 If he said thus, The speckled shall be thy wages;

then all the cattle bare speckled: and if he said thus, The ringstraked shall be thy hire; then bare all the cattle

uariae erunt mercedes tuae pariebant omnes oues uarios fe- ringstraked. tus quando uero e contrario ait alba quaeque accipies pro mercede omnes greges alba pepererunt 31:9 tulitque De- 31:9 Thus God hath taken away the cattle of your father,

and given them to me.

us substantiam patris uestri et dedit mihi 31:10 postquam 31:10 And it came to pass at the time that the cattle con-

ceived, that I lifted up mine eyes, and saw in a dream, and, behold, the rams which leaped upon the cattle were

enim conceptus ouium tempus aduenerat leuaui oculos meos ringstraked, speckled, and grisled. et uidi in somnis ascendentes mares super feminas uarios et maculosos et diuersorum colorum 31:11 dixitque angelus 31:11 And the angel of God spake unto me in a dream,

saying, Jacob: And I said, Here am I.

Dei ad me in somnis Iacob et ego respondi adsum 31:12 31:12 And he said, Lift up now thine eyes, and see, all the rams which leap upon the cattle are ringstraked, speckled, and grisled: for I have seen all that Laban doeth unto thee.

qui ait leua oculos tuos et uide uniuersos masculos ascendentes super feminas uarios respersos atque maculosos uidi enim omnia quae fecit tibi Laban 31:13 ego sum Deus Be- 31:13 I am the God of Bethel, where thou anointedst the

pillar, and where thou vowedst a vow unto me: now arise, get thee out from this land, and return unto the land of thy

thel ubi unxisti lapidem et uotum uouisti mihi nunc ergo kindred.

Biblia Sacra Vulgata

surge et egredere de terra hac reuertens in terram natiuita-

31:14 And Rachel and Leah answered and said unto him, tis tuae 31:14 responderunt Rahel et Lia numquid habemus

Is there yet any portion or inheritance for us in our father’s house?

residui quicquam in facultatibus et hereditate domus patris

31:15 Are we not counted of him strangers? for he hath nostri 31:15 nonne quasi alienas reputauit nos et uendidit

sold us, and hath quite devoured also our money.

31:16 For all the riches which God hath taken from our fa- comeditque pretium nostrum 31:16 sed Deus tulit opes pa-

ther, that is ours, and our children’s: now then, whatsoever God hath said unto thee, do.

tris nostri et nobis eas tradidit ac filiis nostris unde omnia

31:17 Then Jacob rose up, and set his sons and his wives quae praecepit fac 31:17 surrexit itaque Iacob et inpositis

upon camels;

31:18 And he carried away all his cattle, and all his goods liberis et coniugibus suis super camelos abiit 31:18 tulit-

which he had gotten, the cattle of his getting, which he had gotten in Padanaram, for to go to Isaac his father in

the land of Canaan. que omnem substantiam et greges et quicquid in Mesopo-

tamiam quaesierat pergens ad Isaac patrem suum in terram

31:19 And Laban went to shear his sheep: and Rachel had Chanaan 31:19 eo tempore Laban ierat ad tondendas oues et

stolen the images that were her father’s.

31:20 And Jacob stole away unawares to Laban the Syrian, Rahel furata est idola patris sui 31:20 noluitque Iacob con-

in that he told him not that he fled.

31:21 So he fled with all that he had; and he rose up, and fiteri socero quod fugeret 31:21 cumque abisset tam ipse

passed over the river, and set his face toward the mount Gilead.

quam omnia quae iuris eius erant et amne transmisso per-

31:22 And it was told Laban on the third day that Jacob geret contra montem Galaad 31:22 nuntiatum est Laban die

was fled.

31:23 And he took his brethren with him, and pursued af- tertio quod fugeret Iacob 31:23 qui adsumptis fratribus suis

ter him seven days’ journey; and they overtook him in the mount Gilead.

persecutus est eum diebus septem et conprehendit in monte

31:24 And God came to Laban the Syrian in a dream by Galaad 31:24 uiditque in somnis dicentem sibi Dominum

night, and said unto him, Take heed that thou speak not to Jacob either good or bad.

caue ne quicquam aspere loquaris contra Iacob 31:25 iam-

31:25 Then Laban overtook Jacob. Now Jacob had pitched his tent in the mount: and Laban with his brethren pitched in the mount of Gilead.

que Iacob extenderat in monte tabernaculum cum ille consecutus eum cum fratribus suis in eodem monte Galaad fixit

31:26 And Laban said to Jacob, What hast thou done, that tentorium 31:26 et dixit ad Iacob quare ita egisti ut clam

thou hast stolen away unawares to me, and carried away my daughters, as captives taken with the sword?

me abigeres filias meas quasi captiuas gladio 31:27 cur ig-

31:27 Wherefore didst thou flee away secretly, and steal away from me; and didst not tell me, that I might have sent thee away with mirth, and with songs, with tabret, and with harp?

norante me fugere uoluisti nec indicare mihi ut prosequerer

31:28 And hast not suffered me to kiss my sons and my te cum gaudio et canticis et tympanis et cithara 31:28 non

daughters? thou hast now done foolishly in so doing.

es passus ut oscularer filios meos ac filias stulte operatus es

31:29 It is in the power of my hand to do you hurt: but et nunc 31:29 ualet quidem manus mea reddere tibi malum

the God of your father spake unto me yesternight, saying, Take thou heed that thou speak not to Jacob either good or

bad. sed Deus patris uestri heri dixit mihi caue ne loquaris cum 31:30 And now, though thou wouldest needs be gone, Iacob quicquam durius 31:30 esto ad tuos ire cupiebas et

because thou sore longedst after thy father’s house, yet wherefore hast thou stolen my gods?

desiderio tibi erat domus patris tui cur furatus es deos meos

31:31 And Jacob answered and said to Laban, Because I 31:31 respondit Iacob quod inscio te profectus sum timui ne

was afraid: for I said, Peradventure thou wouldest take by force thy daughters from me.

uiolenter auferres filias tuas 31:32 quod autem furti arguis

31:32 With whomsoever thou findest thy gods, let him not live: before our brethren discern thou what is thine with me, and take it to thee. For Jacob knew not that Rachel had stolen them.

apud quemcumque inueneris deos tuos necetur coram fra-

Liber Genesis

tribus nostris scrutare quicquid tuorum apud me inueneris et aufer haec dicens ignorabat quod Rahel furata esset idola 31:33 ingressus itaque Laban tabernaculum Iacob et Liae et 31:33 And Laban went into Jacob’s tent, and into Leah’s

tent, and into the two maidservants’ tents; but he found them not. Then went he out of Leah’s tent, and entered

utriusque famulae non inuenit cumque intrasset tentorium into Rachel’s tent. Rahelis 31:34 illa festinans abscondit idola subter stramen 31:34 Now Rachel had taken the images, and put them

in the camel’s furniture, and sat upon them. And Laban searched all the tent, but found them not.

cameli et sedit desuper scrutantique omne tentorium et nihil inuenienti 31:35 ait ne irascatur dominus meus quod coram 31:35 And she said to her father, Let it not displease my

lord that I cannot rise up before thee; for the custom of women is upon me. And he searched but found not the

te adsurgere nequeo quia iuxta consuetudinem feminarum images. nunc accidit mihi sic delusa sollicitudo quaerentis est 31:36 31:36 And Jacob was wroth, and chode with Laban: and

Jacob answered and said to Laban, What is my trespass? what is my sin, that thou hast so hotly pursued after me?

tumensque Iacob cum iurgio ait quam ob culpam meam et ob quod peccatum sic exarsisti post me 31:37 et scrutatus 31:37 Whereas thou hast searched all my stuff, what hast

thou found of all thy household stuff? set it here before my brethren and thy brethren, that they may judge betwixt

es omnem supellectilem meam quid inuenisti de cuncta sub- us both. stantia domus tuae pone hic coram fratribus meis et fratribus tuis et iudicent inter me et te 31:38 idcirco uiginti annis fui 31:38 This twenty years have I been with thee; thy ewes

and thy she goats have not cast their young, and the rams of thy flock have I not eaten.

tecum oues tuae et caprae steriles non fuerunt arietes gregis tui non comedi 31:39 nec captum a bestia ostendi tibi 31:39 That which was torn of beasts I brought not unto

thee; I bare the loss of it; of my hand didst thou require it, whether stolen by day, or stolen by night.

ego damnum omne reddebam quicquid furto perierat a me exigebas 31:40 die noctuque aestu urebar et gelu fugiebat 31:40 Thus I was; in the day the drought consumed me,

and the frost by night; and my sleep departed from mine eyes.

somnus ab oculis meis 31:41 sic per uiginti annos in domo

31:41 Thus have I been twenty years in thy house; I served thee fourteen years for thy two daughters, and six years for thy cattle: and thou hast changed my wages ten times.

tua seruiui tibi quattuordecim pro filiabus et sex pro gregibus tuis inmutasti quoque mercedem meam decem uicibus 31:42 nisi Deus patris mei Abraham et Timor Isaac adfuis- 31:42 Except the God of my father, the God of Abraham,

and the fear of Isaac, had been with me, surely thou hadst sent me away now empty. God hath seen mine affliction

set mihi forsitan modo nudum me dimisisses adflictionem and the labour of my hands, and rebuked thee yesternight.

meam et laborem manuum mearum respexit Deus et arguit te heri 31:43 respondit ei Laban filiae et filii et greges tui 31:43 And Laban answered and said unto Jacob, These

daughters are my daughters, and these children are my children, and these cattle are my cattle, and all that thou seest is mine: and what can I do this day unto these my daughters, or unto their children which they have born?

et omnia quae cernis mea sunt quid possum facere filiis et nepotibus meis 31:44 ueni ergo et ineamus foedus ut sit te- 31:44 Now therefore come thou, let us make a covenant, I

and thou; and let it be for a witness between me and thee.

stimonium inter me et te 31:45 tulit itaque Iacob lapidem et 31:45 And Jacob took a stone, and set it up for a pillar. erexit illum in titulum 31:46 dixitque fratribus suis adferte 31:46 And Jacob said unto his brethren, Gather stones; and

they took stones, and made an heap: and they did eat there upon the heap.

lapides qui congregantes fecerunt tumulum comederuntque super eum 31:47 quem uocauit Laban tumulus Testis et Ia- 31:47 And Laban called it Jegarsahadutha: but Jacob

called it Galeed.

cob aceruum Testimonii uterque iuxta proprietatem linguae suae 31:48 dixitque Laban tumulus iste testis erit inter me 31:48 And Laban said, This heap is a witness between me

and thee this day. Therefore was the name of it called Galeed;

Biblia Sacra Vulgata

et te hodie et idcirco appellatum est nomen eius Galaad id

31:49 And Mizpah; for he said, The LORD watch between est tumulus Testis 31:49 intueatur Dominus et iudicet inter

me and thee, when we are absent one from another.

31:50 If thou shalt afflict my daughters, or if thou shalt nos quando recesserimus a nobis 31:50 si adflixeris filias

take other wives beside my daughters, no man is with us; see, God is witness betwixt me and thee.

meas et si introduxeris uxores alias super eas nullus sermo-

31:51 And Laban said to Jacob, Behold this heap, and be- nis nostri testis est absque Deo qui praesens respicit 31:51

hold this pillar, which I have cast betwixt me and thee:

dixitque rursus ad Iacob en tumulus hic et lapis quem erexi

31:52 This heap be witness, and this pillar be witness, that inter me et te 31:52 testis erit tumulus inquam iste et lapis I will not pass over this heap to thee, and that thou shalt not pass over this heap and this pillar unto me, for harm.

sint in testimonio si aut ego transiero illum pergens ad te

31:53 The God of Abraham, and the God of Nahor, the aut tu praeterieris malum mihi cogitans 31:53 Deus Abra-

God of their father, judge betwixt us. And Jacob sware by the fear of his father Isaac.

ham et Deus Nahor iudicet inter nos Deus patris eorum iura-

31:54 Then Jacob offered sacrifice upon the mount, and uit Iacob per Timorem patris sui Isaac 31:54 immolatisque

called his brethren to eat bread: and they did eat bread, and tarried all night in the mount.

uictimis in monte uocauit fratres suos ut ederent panem qui

31:55 And early in the morning Laban rose up, and kissed cum comedissent manserunt ibi 31:55 Laban uero de nocte

his sons and his daughters, and blessed them: and Laban departed, and returned unto his place.

consurgens osculatus est filios et filias suas et benedixit illis reuersus in locum suum

32:1 And Jacob went on his way, and the angels of God 32:1 Iacob quoque abiit itinere quo coeperat fueruntque

met him.

32:2 And when Jacob saw them, he said, This is God’s ei obuiam angeli Dei 32:2 quos cum uidisset ait castra Dei

host: and he called the name of that place Mahanaim.

sunt haec et appellauit nomen loci illius Manaim id est Cas-

32:3 And Jacob sent messengers before him to Esau his tra 32:3 misit autem et nuntios ante se ad Esau fratrem

brother unto the land of Seir, the country of Edom.

32:4 And he commanded them, saying, Thus shall ye suum in terram Seir regionis Edom 32:4 praecepitque eis

speak unto my lord Esau; Thy servant Jacob saith thus, I have sojourned with Laban, and stayed there until now:

dicens sic loquimini domino meo Esau haec dicit frater tuus Iacob apud Laban peregrinatus sum et fui usque in praesen-

32:5 And I have oxen, and asses, flocks, and menservants, tem diem 32:5 habeo boues et asinos oues et seruos atque and womenservants: and I have sent to tell my lord, that I may find grace in thy sight.

ancillas mittoque nunc legationem ad dominum meum ut in-

32:6 And the messengers returned to Jacob, saying, We ueniam gratiam in conspectu tuo 32:6 reuersi sunt nuntii ad

came to thy brother Esau, and also he cometh to meet thee, and four hundred men with him.

Iacob dicentes uenimus ad Esau fratrem tuum et ecce pro-

32:7 Then Jacob was greatly afraid and distressed: and he perat in occursum tibi cum quadringentis uiris 32:7 timuit

divided the people that was with him, and the flocks, and herds, and the camels, into two bands;

Iacob ualde et perterritus diuisit populum qui secum erat gre-

32:8 And said, If Esau come to the one company, and ges quoque et oues et boues et camelos in duas turmas 32:8

smite it, then the other company which is left shall escape.

dicens si uenerit Esau ad unam turmam et percusserit eam

32:9 And Jacob said, O God of my father Abraham, and alia turma quae reliqua est saluabitur 32:9 dixitque Iacob

God of my father Isaac, the LORD which saidst unto me, Return unto thy country, and to thy kindred, and I will deal

well with thee: Deus patris mei Abraham et Deus patris mei Isaac Domine qui dixisti mihi reuertere in terram tuam et in locum natiui-

Liber Genesis

tatis tuae et benefaciam tibi 32:10 minor sum cunctis mise- 32:10 I am not worthy of the least of all the mercies, and

of all the truth, which thou hast shewed unto thy servant; for with my staff I passed over this Jordan; and now I am

rationibus et ueritate quam explesti seruo tuo in baculo meo become two bands. transiui Iordanem istum et nunc cum duabus turmis regredior 32:11 erue me de manu fratris mei de manu Esau quia 32:11 Deliver me, I pray thee, from the hand of my

brother, from the hand of Esau: for I fear him, lest he will come and smite me, and the mother with the children.

ualde eum timeo ne forte ueniens percutiat matrem cum filiis 32:12 tu locutus es quod bene mihi faceres et dilatares semen 32:12 And thou saidst, I will surely do thee good, and

make thy seed as the sand of the sea, which cannot be numbered for multitude.

meum sicut harenam maris quae prae multitudine numerari non potest 32:13 cumque dormisset ibi nocte illa separauit 32:13 And he lodged there that same night; and took of

that which came to his hand a present for Esau his brother;

de his quae habebat munera Esau fratri suo 32:14 capras 32:14 Two hundred she goats, and twenty he goats, two

hundred ewes, and twenty rams,

ducentas hircos uiginti oues ducentas arietes uiginti 32:15 32:15 Thirty milch camels with their colts, forty kine, and

ten bulls, twenty she asses, and ten foals.

camelos fetas cum pullis suis triginta uaccas quadraginta et tauros uiginti asinas uiginti et pullos earum decem 32:16 32:16 And he delivered them into the hand of his servants,

every drove by themselves; and said unto his servants, Pass over before me, and put a space betwixt drove and

et misit per manus seruorum suorum singulos seorsum gre- drove. ges dixitque pueris suis antecedite me et sit spatium inter gregem et gregem 32:17 et praecepit priori dicens si ob- 32:17 And he commanded the foremost, saying, When

Esau my brother meeteth thee, and asketh thee, saying, Whose art thou? and whither goest thou? and whose are

uium habueris Esau fratrem meum et interrogauerit te cuius these before thee?

es et quo uadis et cuius sunt ista quae sequeris 32:18 re- 32:18 Then thou shalt say, They be thy servant Jacob’s; it is a present sent unto my lord Esau: and, behold, also he is behind us.

spondebis serui tui Iacob munera misit domino meo Esau ipse quoque post nos uenit 32:19 similiter mandata dedit se- 32:19 And so commanded he the second, and the third,

and all that followed the droves, saying, On this manner shall ye speak unto Esau, when ye find him.

cundo ac tertio et cunctis qui sequebantur greges dicens hisdem uerbis loquimini ad Esau cum inueneritis eum 32:20 32:20 And say ye moreover, Behold, thy servant Jacob is

behind us. For he said, I will appease him with the present that goeth before me, and afterward I will see his face;

et addetis ipse quoque seruus tuus Iacob iter nostrum inse- peradventure he will accept of me. quitur dixit enim placabo illum muneribus quae praecedunt et postea uidebo forsitan propitiabitur mihi 32:21 praeces- 32:21 So went the present over before him: and himself

lodged that night in the company.

serunt itaque munera ante eum ipse uero mansit nocte illa in Castris 32:22 cumque mature surrexisset tulit duas uxo- 32:22 And he rose up that night, and took his two wives,

and his two womenservants, and his eleven sons, and passed over the ford Jabbok.

res suas et totidem famulas cum undecim filiis et transiuit uadum Iaboc 32:23 transductisque omnibus quae ad se per- 32:23 And he took them, and sent them over the brook,

and sent over that he had.

tinebant 32:24 remansit solus et ecce uir luctabatur cum eo 32:24 And Jacob was left alone; and there wrestled a man

with him until the breaking of the day.

usque mane 32:25 qui cum uideret quod eum superare non 32:25 And when he saw that he prevailed not against him, he touched the hollow of his thigh; and the hollow of Jacob’s thigh was out of joint, as he wrestled with him.

posset tetigit neruum femoris eius et statim emarcuit 32:26

32:26 And he said, Let me go, for the day breaketh. And he said, I will not let thee go, except thou bless me.

dixitque ad eum dimitte me iam enim ascendit aurora respondit non dimittam te nisi benedixeris mihi 32:27 ait ergo 32:27 And he said unto him, What is thy name? And he

said, Jacob.

Biblia Sacra Vulgata

32:28 And he said, Thy name shall be called no more Ja- quod nomen est tibi respondit Iacob 32:28 at ille nequa-

cob, but Israel: for as a prince hast thou power with God and with men, and hast prevailed.

quam inquit Iacob appellabitur nomen tuum sed Israhel quoniam si contra Deum fortis fuisti quanto magis contra homi-

32:29 And Jacob asked him, and said, Tell me, I pray thee, nes praeualebis 32:29 interrogauit eum Iacob dic mihi quo

thy name. And he said, Wherefore is it that thou dost ask after my name? And he blessed him there.

appellaris nomine respondit cur quaeris nomen meum et be-

32:30 And Jacob called the name of the place Peniel: for nedixit ei in eodem loco 32:30 uocauitque Iacob nomen loci

I have seen God face to face, and my life is preserved.

illius Phanuhel dicens uidi Deum facie ad faciem et salua

32:31 And as he passed over Penuel the sun rose upon him, facta est anima mea 32:31 ortusque est ei statim sol post-

and he halted upon his thigh.

quam transgressus est Phanuhel ipse uero claudicabat pede

32:32 Therefore the children of Israel eat not of the sinew 32:32 quam ob causam non comedunt filii Israhel neruum qui which shrank, which is upon the hollow of the thigh, unto this day: because he touched the hollow of Jacob’s thigh

in the sinew that shrank. emarcuit in femore Iacob usque in praesentem diem eo quod tetigerit neruum femoris eius et obstipuerit

33:1 And Jacob lifted up his eyes, and looked, and, be- 33:1 leuans autem Iacob oculos suos uidit uenientem Esau

hold, Esau came, and with him four hundred men. And he divided the children unto Leah, and unto Rachel, and unto

the two handmaids. et cum eo quadringentos uiros diuisitque filios Liae et Ra-33:2 And he put the handmaids and their children fore- hel ambarumque famularum 33:2 et posuit utramque ancil-

most, and Leah and her children after, and Rachel and Joseph hindermost.

lam et liberos earum in principio Liam uero et filios eius

33:3 And he passed over before them, and bowed him- in secundo loco Rahel autem et Ioseph nouissimos 33:3 et

self to the ground seven times, until he came near to his brother.

ipse praegrediens adorauit pronus in terram septies donec

33:4 And Esau ran to meet him, and embraced him, and adpropinquaret frater eius 33:4 currens itaque Esau obuiam

fell on his neck, and kissed him: and they wept.

fratri suo amplexatus est eum stringensque collum et oscu-

33:5 And he lifted up his eyes, and saw the women and the lans fleuit 33:5 leuatisque oculis uidit mulieres et paruulos

children; and said, Who are those with thee? And he said, The children which God hath graciously given thy servant.

earum et ait quid sibi uolunt isti et si ad te pertinent respon-

33:6 Then the handmaidens came near, they and their chil- dit paruuli sunt quos donauit mihi Deus seruo tuo 33:6 et dren, and they bowed themselves.

33:7 And Leah also with her children came near, and adpropinquantes ancillae et filii earum incuruati sunt 33:7

bowed themselves: and after came Joseph near and Rachel, and they bowed themselves.

accessitque Lia cum liberis suis et cum similiter adorassent

33:8 And he said, What meanest thou by all this drove extremi Ioseph et Rahel adorauerunt 33:8 quaenam sunt in-

which I met? And he said, These are to find grace in the sight of my lord.

quit istae turmae quas obuias habui respondit ut inuenirem

33:9 And Esau said, I have enough, my brother; keep that gratiam coram domino meo 33:9 et ille habeo ait plurima thou hast unto thyself.

33:10 And Jacob said, Nay, I pray thee, if now I have frater mi sint tua tibi 33:10 dixit Iacob noli ita obsecro sed

found grace in thy sight, then receive my present at my hand: for therefore I have seen thy face, as though I had

seen the face of God, and thou wast pleased with me. si inueni gratiam in oculis tuis accipe munusculum de ma-

nibus meis sic enim uidi faciem tuam quasi uiderim uul-

33:11 Take, I pray thee, my blessing that is brought to tum Dei esto mihi propitius 33:11 et suscipe benedictionem

thee; because God hath dealt graciously with me, and because I have enough. And he urged him, and he took it.

quam adtuli tibi et quam donauit mihi Deus tribuens omnia

Liber Genesis

uix fratre conpellente suscipiens 33:12 ait gradiamur simul 33:12 And he said, Let us take our journey, and let us go,

and I will go before thee.

eroque socius itineris tui 33:13 dixit Iacob nosti domine mi 33:13 And he said unto him, My lord knoweth that the

children are tender, and the flocks and herds with young are with me: and if men should overdrive them one day,

quod paruulos habeam teneros et oues ac boues fetas me- all the flock will die. cum quas si plus in ambulando fecero laborare morientur una die cuncti greges 33:14 praecedat dominus meus ante 33:14 Let my lord, I pray thee, pass over before his ser-

vant: and I will lead on softly, according as the cattle that goeth before me and the children be able to endure, until I

seruum suum et ego sequar paulatim uestigia eius sicut ui- come unto my lord unto Seir. dero posse paruulos meos donec ueniam ad dominum meum in Seir 33:15 respondit Esau oro te ut de populo qui mecum 33:15 And Esau said, Let me now leave with thee some of

the folk that are with me. And he said, What needeth it? let me find grace in the sight of my lord.

est saltem socii remaneant uiae tuae non est inquit necesse hoc uno indigeo ut inueniam gratiam in conspectu domini mei 33:16 reuersus est itaque illo die Esau itinere quo ue- 33:16 So Esau returned that day on his way unto Seir. nerat in Seir 33:17 et Iacob uenit in Soccoth ubi aedificata 33:17 And Jacob journeyed to Succoth, and built him an

house, and made booths for his cattle: therefore the name of the place is called Succoth.

domo et fixis tentoriis appellauit nomen loci illius Soccoth id est Tabernacula 33:18 transiuitque in Salem urbem Sy- 33:18 And Jacob came to Shalem, a city of Shechem,

which is in the land of Canaan, when he came from Padanaram; and pitched his tent before the city.

cimorum quae est in terra Chanaan postquam regressus est de Mesopotamiam Syriae et habitauit iuxta oppidum 33:19 33:19 And he bought a parcel of a field, where he had

spread his tent, at the hand of the children of Hamor, Shechem’s father, for an hundred pieces of money.

emitque partem agri in qua fixerat tabernaculum a filiis Emor patris Sychem centum agnis 33:20 et erecto ibi altari inuo- 33:20 And he erected there an altar, and called it Elelo-

heIsrael.

cauit super illud Fortissimum Deum Israhel

34:1 egressa est autem Dina filia Liae ut uideret mulie- 34:1 And Dinah the daughter of Leah, which she bare unto

Jacob, went out to see the daughters of the land.

res regionis illius 34:2 quam cum uidisset Sychem filius 34:2 And when Shechem the son of Hamor the Hivite,

prince of the country, saw her, he took her, and lay with her, and defiled her.

Emor Euei princeps terrae illius adamauit et rapuit et dormiuit cum illa ui opprimens uirginem 34:3 et conglutinata 34:3 And his soul clave unto Dinah the daughter of Ja-

cob, and he loved the damsel, and spake kindly unto the damsel.

est anima eius cum ea tristemque blanditiis deliniuit 34:4

34:4 And Shechem spake unto his father Hamor, saying, Get me this damsel to wife.

et pergens ad Emor patrem suum accipe mihi inquit puellam hanc coniugem 34:5 quod cum audisset Iacob absenti- 34:5 And Jacob heard that he had defiled Dinah his daugh-

ter: now his sons were with his cattle in the field: and Jacob held his peace until they were come.

bus filiis et in pastu occupatis pecorum siluit donec rediient 34:6 egresso autem Emor patre Sychem ut loqueretur ad Ia- 34:6 And Hamor the father of Shechem went out unto Ja-

cob to commune with him.

cob 34:7 ecce filii eius ueniebant de agro auditoque quod 34:7 And the sons of Jacob came out of the field when

they heard it: and the men were grieved, and they were very wroth, because he had wrought folly in Israel in lying

acciderat irati sunt ualde eo quod foedam rem esset opera- with Jacob’s daughter: which thing ought not to be done. tus in Israhel et uiolata filia Iacob rem inlicitam perpetrasset 34:8 locutus est itaque Emor ad eos Sychem filii mei adhesit 34:8 And Hamor communed with them, saying, The soul

of my son Shechem longeth for your daughter: I pray you give her him to wife.

anima filiae uestrae date eam illi uxorem 34:9 et iungamus

34:9 And make ye marriages with us, and give your daughters unto us, and take our daughters unto you.

Biblia Sacra Vulgata

uicissim conubia filias uestras tradite nobis et filias nostras

34:10 And ye shall dwell with us: and the land shall be accipite 34:10 et habitate nobiscum terra in potestate ues-

before you; dwell and trade ye therein, and get you possessions therein.

tra est exercete negotiamini et possidete eam 34:11 sed et

34:11 And Shechem said unto her father and unto her brethren, Let me find grace in your eyes, and what ye shall say unto me I will give.

Sychem ad patrem et ad fratres eius ait inueniam gratiam

34:12 Ask me never so much dowry and gift, and I will coram uobis et quaecumque statueritis dabo 34:12 augete

give according as ye shall say unto me: but give me the damsel to wife.

dotem munera postulate libens tribuam quod petieritis tan-

34:13 And the sons of Jacob answered Shechem and tum date mihi puellam hanc uxorem 34:13 responderunt fi-

Hamor his father deceitfully, and said, because he had defiled Dinah their sister:

lii Iacob Sychem et patri eius in dolo saeuientes ob stuprum

34:14 And they said unto them, We cannot do this thing, to sororis 34:14 non possumus facere quod petitis nec dare so-

give our sister to one that is uncircumcised; for that were a reproach unto us:

rorem nostram homini incircumciso quod inlicitum et nefa-

34:15 But in this will we consent unto you: If ye will be rium est apud nos 34:15 sed in hoc ualebimus foederari si

as we be, that every male of you be circumcised;

esse uolueritis nostri similes et circumcidatur in uobis omne

34:16 Then will we give our daughters unto you, and we masculini sexus 34:16 tunc dabimus et accipiemus mutuo

will take your daughters to us, and we will dwell with you, and we will become one people.

filias nostras ac uestras et habitabimus uobiscum erimusque

34:17 But if ye will not hearken unto us, to be circumcised; unus populus 34:17 sin autem circumcidi nolueritis tolle-

then will we take our daughter, and we will be gone.

34:18 And their words pleased Hamor, and Shechem mus filiam nostram et recedemus 34:18 placuit oblatio eo-

Hamor’s son.

34:19 And the young man deferred not to do the thing, rum Emor et Sychem filio eius 34:19 nec distulit adulescens

because he had delight in Jacob’s daughter: and he was more honourable than all the house of his father.

quin statim quod petebatur expleret amabat enim puellam

34:20 And Hamor and Shechem his son came unto the ualde et ipse erat inclitus in omni domo patris sui 34:20

gate of their city, and communed with the men of their city, saying,

ingressique portam urbis locuti sunt populo 34:21 uiri isti

34:21 These men are peaceable with us; therefore let them dwell in the land, and trade therein; for the land, behold, it is large enough for them; let us take their daughters to us for wives, and let us give them our daughters.

pacifici sunt et uolunt habitare nobiscum negotientur in terra et exerceant eam quae spatiosa et lata cultoribus indiget fi-

34:22 Only herein will the men consent unto us for to lias eorum accipiemus uxores et nostras illis dabimus 34:22

dwell with us, to be one people, if every male among us be circumcised, as they are circumcised.

unum est quod differtur tantum bonum si circumcidamus

34:23 Shall not their cattle and their substance and every masculos nostros ritum gentis imitantes 34:23 et substantia

beast of their’s be our’s? only let us consent unto them, and they will dwell with us.

eorum et pecora et cuncta quae possident nostra erunt tantum in hoc adquiescamus et habitantes simul unum efficie-

34:24 And unto Hamor and unto Shechem his son hear- mus populum 34:24 adsensi sunt omnes circumcisis cunctis

kened all that went out of the gate of his city; and every male was circumcised, all that went out of the gate of his

city. maribus 34:25 et ecce die tertio quando grauissimus uulne-

34:25 And it came to pass on the third day, when they were sore, that two of the sons of Jacob, Simeon and Levi, Dinah’s brethren, took each man his sword, and came upon the city boldly, and slew all the males.

rum dolor est arreptis duo Iacob filii Symeon et Leui fratres Dinae gladiis ingressi sunt urbem confidenter interfectisque

34:26 And they slew Hamor and Shechem his son with omnibus masculis 34:26 Emor et Sychem pariter necaue-

the edge of the sword, and took Dinah out of Shechem’s house, and went out.

runt tollentes Dinam de domo Sychem sororem suam 34:27

34:27 The sons of Jacob came upon the slain, and spoiled the city, because they had defiled their sister.

Liber Genesis

quibus egressis inruerunt super occisos ceteri filii Iacob et depopulati sunt urbem in ultionem stupri 34:28 oues eorum 34:28 They took their sheep, and their oxen, and their

asses, and that which was in the city, and that which was in the field,

et armenta et asinos cunctaque uastantes quae in domibus

et in agris erant 34:29 paruulos quoque et uxores eorum 34:29 And all their wealth, and all their little ones, and their wives took they captive, and spoiled even all that was in the house.

duxere captiuas 34:30 quibus patratis audacter Iacob dixit

34:30 And Jacob said to Simeon and Levi, Ye have troubled me to make me to stink among the inhabitants of the land, among the Canaanites and the Perizzites: and I being few in number, they shall gather themselves together against me, and slay me; and I shall be destroyed, I and my house.

ad Symeon et Leui turbastis me et odiosum fecistis Chananeis et Ferezeis habitatoribus terrae huius nos pauci sumus illi congregati percutient me et delebor ego et domus mea 34:31 responderunt numquid ut scorto abuti debuere sorore 34:31 And they said, Should he deal with our sister as with

an harlot?

nostra

35:1 interea locutus est Deus ad Iacob surge et ascende Be- 35:1 And God said unto Jacob, Arise, go up to Bethel,

and dwell there: and make there an altar unto God, that appeared unto thee when thou fleddest from the face of

thel et habita ibi facque altare Deo qui apparuit tibi quando Esau thy brother.

fugiebas Esau fratrem tuum 35:2 Iacob uero conuocata 35:2 Then Jacob said unto his household, and to all that were with him, Put away the strange gods that are among you, and be clean, and change your garments:

omni domo sua ait abicite deos alienos qui in medio uestri sunt et mundamini ac mutate uestimenta uestra 35:3 surgite 35:3 And let us arise, and go up to Bethel; and I will make

there an altar unto God, who answered me in the day of my distress, and was with me in the way which I went.

et ascendamus in Bethel ut faciamus ibi altare Deo qui exaudiuit me in die tribulationis meae et fuit socius itineris mei 35:4 dederunt ergo ei omnes deos alienos quos habebant et 35:4 And they gave unto Jacob all the strange gods which

were in their hand, and all their earrings which were in their ears; and Jacob hid them under the oak which was by

inaures quae erant in auribus eorum at ille infodit ea sub- Shechem. ter terebinthum quae est post urbem Sychem 35:5 cumque 35:5 And they journeyed: and the terror of God was upon

the cities that were round about them, and they did not pursue after the sons of Jacob.

profecti essent terror Dei inuasit omnes per circuitum ciuitates et non sunt ausi persequi recedentes 35:6 uenit igitur 35:6 So Jacob came to Luz, which is in the land of Canaan,

that is, Bethel, he and all the people that were with him.

Iacob Luzam quae est in terra Chanaan cognomento Bethel

ipse et omnis populus cum eo 35:7 aedificauitque ibi al- 35:7 And he built there an altar, and called the place Elbethel: because there God appeared unto him, when he fled from the face of his brother.

tare et appellauit nomen loci Domus Dei ibi enim apparuit ei Deus cum fugeret fratrem suum 35:8 eodem tempore mor- 35:8 But Deborah Rebekah’s nurse died, and she was

buried beneath Bethel under an oak: and the name of it was called Allonbachuth.

tua est Debbora nutrix Rebeccae et sepulta ad radices Bethel subter quercum uocatumque est nomen loci quercus Fletus 35:9 apparuit autem iterum Deus Iacob postquam reuersus 35:9 And God appeared unto Jacob again, when he came

out of Padanaram, and blessed him.

est de Mesopotamiam Syriae benedixitque ei 35:10 dicens 35:10 And God said unto him, Thy name is Jacob: thy

name shall not be called any more Jacob, but Israel shall be thy name: and he called his name Israel.

non uocaberis ultra Iacob sed Israhel erit nomen tuum et appellauit eum Israhel 35:11 dixitque ei ego Deus omnipotens 35:11 And God said unto him, I am God Almighty: be

fruitful and multiply; a nation and a company of nations shall be of thee, and kings shall come out of thy loins;

cresce et multiplicare gentes et populi nationum erunt ex te

Biblia Sacra Vulgata

35:12 And the land which I gave Abraham and Isaac, to reges de lumbis tuis egredientur 35:12 terramque quam de-

thee I will give it, and to thy seed after thee will I give the land.

di Abraham et Isaac dabo tibi et semini tuo post te 35:13 et

35:13 And God went up from him in the place where he talked with him.

35:14 And Jacob set up a pillar in the place where he recessit ab eo 35:14 ille uero erexit titulum lapideum in loco

talked with him, even a pillar of stone: and he poured a drink offering thereon, and he poured oil thereon.

quo locutus ei fuerat Deus libans super eum libamina et ef-

35:15 And Jacob called the name of the place where God fundens oleum 35:15 uocansque nomen loci Bethel 35:16

spake with him, Bethel. 35:16 And they journeyed from Bethel; and there was but a little way to come to Ephrath: and Rachel travailed, and she had hard labour.

egressus inde uenit uerno tempore ad terram quae ducit Ef-

35:17 And it came to pass, when she was in hard labour, ratham in qua cum parturiret Rahel 35:17 ob difficultatem

that the midwife said unto her, Fear not; thou shalt have this son also.

partus periclitari coepit dixitque ei obsetrix noli timere quia

35:18 And it came to pass, as her soul was in departing, et hunc habebis filium 35:18 egrediente autem anima prae

(for she died) that she called his name Benoni: but his father called him Benjamin.

dolore et inminente iam morte uocauit nomen filii sui Benoni id est filius doloris mei pater uero appellauit eum Be-

35:19 And Rachel died, and was buried in the way to niamin id est filius dexterae 35:19 mortua est ergo Rahel et

Ephrath, which is Bethlehem.

35:20 And Jacob set a pillar upon her grave: that is the sepulta in uia quae ducit Efratham haec est Bethleem 35:20

pillar of Rachel’s grave unto this day.

erexitque Iacob titulum super sepulchrum eius hic est titulus

35:21 And Israel journeyed, and spread his tent beyond monumenti Rahel usque in praesentem diem 35:21 egressus

the tower of Edar.

35:22 And it came to pass, when Israel dwelt in that land, inde fixit tabernaculum trans turrem Gregis 35:22 cumque

that Reuben went and lay with Bilhah his father’s concubine: and Israel heard it. Now the sons of Jacob were

twelve: habitaret in illa regione abiit Ruben et dormiuit cum Bala

concubina patris sui quod illum minime latuit erant autem

35:23 The sons of Leah; Reuben, Jacob’s firstborn, and filii Iacob duodecim 35:23 filii Liae primogenitus Ruben et

Simeon, and Levi, and Judah, and Issachar, and Zebulun:

35:24 The sons of Rachel; Joseph, and Benjamin: Symeon et Leui et Iudas et Isachar et Zabulon 35:24 filii 35:25 And the sons of Bilhah, Rachel’s handmaid; Dan, Rahel Ioseph et Beniamin 35:25 filii Balae ancillae Rahelis

and Naphtali:

35:26 And the sons of Zilpah, Leah’s handmaid: Gad, and Dan et Nepthalim 35:26 filii Zelphae ancillae Liae Gad et

Asher: these are the sons of Jacob, which were born to him in Padanaram.

Aser hii filii Iacob qui nati sunt ei in Mesopotamiam Syriae

35:27 And Jacob came unto Isaac his father unto Mamre, 35:27 uenit etiam ad Isaac patrem suum in Mambre ciuita-

unto the city of Arbah, which is Hebron, where Abraham and Isaac sojourned.

tem Arbee haec est Hebron in qua peregrinatus est Abraham

35:28 And the days of Isaac were an hundred and et Isaac 35:28 et conpleti sunt dies Isaac centum octoginta

fourscore years.

35:29 And Isaac gave up the ghost, and died, and was gath- annorum 35:29 consumptusque aetate mortuus est et adpo-

ered unto his people, being old and full of days: and his sons Esau and Jacob buried him.

situs populo suo senex et plenus dierum et sepelierunt eum Esau et Iacob filii sui

36:1 Now these are the generations of Esau, who is Edom. 36:1 hae sunt autem generationes Esau ipse est Edom

36:2 Esau took his wives of the daughters of Canaan; 36:2 Esau accepit uxores de filiabus Chanaan Ada filiam

Adah the daughter of Elon the Hittite, and Aholibamah the daughter of Anah the daughter of Zibeon the Hivite;

Elom Hetthei et Oolibama filiam Anae filiae Sebeon Euei

36:3 And Bashemath Ishmael’s daughter, sister of Neba- 36:3 Basemath quoque filiam Ismahel sororem Nabaioth joth.

Liber Genesis

36:4 peperit autem Ada Eliphaz Basemath genuit Rauhel 36:4 And Adah bare to Esau Eliphaz; and Bashemath bare

Reuel;

36:5 Oolibama edidit Hieus et Hielom et Core hii filii Esau 36:5 And Aholibamah bare Jeush, and Jaalam, and Korah:

these are the sons of Esau, which were born unto him in the land of Canaan.

qui nati sunt ei in terra Chanaan 36:6 tulit autem Esau uxo-

36:6 And Esau took his wives, and his sons, and his daughters, and all the persons of his house, and his cattle, and all his beasts, and all his substance, which he had got in the land of Canaan; and went into the country from the face of his brother Jacob.

res suas et filios et filias et omnem animam domus suae et substantiam et pecora et cuncta quae habere poterat in terra Chanaan et abiit in alteram regionem recessitque a fratre suo Iacob 36:7 diuites enim erant ualde et simul habitare non 36:7 For their riches were more than that they might dwell

together; and the land wherein they were strangers could not bear them because of their cattle.

poterant nec sustinebat eos terra peregrinationis eorum prae multitudine gregum 36:8 habitauitque Esau in monte Seir 36:8 Thus dwelt Esau in mount Seir: Esau is Edom. ipse est Edom 36:9 hae sunt generationes Esau patris Edom 36:9 And these are the generations of Esau the father of

the Edomites in mount Seir:

in monte Seir 36:10 et haec nomina filiorum eius Eliphaz 36:10 These are the names of Esau’s sons; Eliphaz the son of Adah the wife of Esau, Reuel the son of Bashemath the wife of Esau.

filius Ada uxoris Esau Rauhel quoque filius Basemath uxoris eius 36:11 fueruntque filii Eliphaz Theman Omar Se- 36:11 And the sons of Eliphaz were Teman, Omar, Zepho,

and Gatam, and Kenaz.

phu et Gatham et Cenez 36:12 erat autem Thamna concu- 36:12 And Timna was concubine to Eliphaz Esau’s son;

and she bare to Eliphaz Amalek: these were the sons of Adah Esau’s wife.

bina Eliphaz filii Esau quae peperit ei Amalech hii sunt filii

Adae uxoris Esau 36:13 filii autem Rauhel Naath et Zara 36:13 And these are the sons of Reuel; Nahath, and Zerah, Shammah, and Mizzah: these were the sons of Bashemath Esau’s wife.

Semma et Meza hii filii Basemath uxoris Esau 36:14 isti

36:14 And these were the sons of Aholibamah, the daughter of Anah the daughter of Zibeon, Esau’s wife: and she bare to Esau Jeush, and Jaalam, and Korah.

quoque erant filii Oolibama filiae Ana filiae Sebeon uxoris Esau quos genuit ei Hieus et Hielom et Core 36:15 hii du- 36:15 These were dukes of the sons of Esau: the sons

of Eliphaz the firstborn son of Esau; duke Teman, duke Omar, duke Zepho, duke Kenaz,

ces filiorum Esau filii Eliphaz primogeniti Esau dux Theman

dux Omar dux Sephu dux Cenez 36:16 dux Core dux Gat- 36:16 Duke Korah, duke Gatam, and duke Amalek: these are the dukes that came of Eliphaz in the land of Edom; these were the sons of Adah.

ham dux Amalech hii filii Eliphaz in terra Edom et hii filii Adae 36:17 hii quoque filii Rauhel filii Esau dux Naath dux 36:17 And these are the sons of Reuel Esau’s son; duke

Nahath, duke Zerah, duke Shammah, duke Mizzah: these are the dukes that came of Reuel in the land of Edom;

Zara dux Semma dux Meza hii duces Rauhel in terra Edom these are the sons of Bashemath Esau’s wife. isti filii Basemath uxoris Esau 36:18 hii autem filii Ooli- 36:18 And these are the sons of Aholibamah Esau’s wife;

duke Jeush, duke Jaalam, duke Korah: these were the dukes that came of Aholibamah the daughter of Anah,

bama uxoris Esau dux Hieus dux Hielom dux Core hii duces Esau’s wife. Oolibama filiae Ana uxoris Esau 36:19 isti filii Esau et hii 36:19 These are the sons of Esau, who is Edom, and these

are their dukes.

duces eorum ipse est Edom 36:20 isti filii Seir Horrei habi- 36:20 These are the sons of Seir the Horite, who inhabited

the land; Lotan, and Shobal, and Zibeon, and Anah,

tatores terrae Lotham et Sobal et Sebeon et Anan 36:21 Di- 36:21 And Dishon, and Ezer, and Dishan: these are the

dukes of the Horites, the children of Seir in the land of Edom.

son et Eser et Disan hii duces Horrei filii Seir in terra Edom 36:22 facti sunt autem filii Lotham Horrei et Heman erat au- 36:22 And the children of Lotan were Hori and Hemam;

and Lotan’s sister was Timna.

tem soror Lotham Thamna 36:23 et isti filii Sobal Aluam et 36:23 And the children of Shobal were these; Alvan, and

Manahath, and Ebal, Shepho, and Onam.

Maneeth et Hebal Sephi et Onam 36:24 et hii filii Sebeon 36:24 And these are the children of Zibeon; both Ajah,

and Anah: this was that Anah that found the mules in the wilderness, as he fed the asses of Zibeon his father.

Biblia Sacra Vulgata

Ahaia et Anam iste est Ana qui inuenit aquas calidas in soli-

36:25 And the children of Anah were these; Dishon, and tudine cum pasceret asinos Sebeon patris sui 36:25 habuit-

Aholibamah the daughter of Anah.

36:26 And these are the children of Dishon; Hemdan, and que filium Disan et filiam Oolibama 36:26 et isti filii Disan

Eshban, and Ithran, and Cheran. 36:27 The children of Ezer are these; Bilhan, and Zaavan, Amdan et Esban et Iethran et Charan 36:27 hii quoque filii and Akan.

36:28 The children of Dishan are these; Uz, and Aran. Eser Balaan et Zeuan et Acham 36:28 habuit autem filios

36:29 These are the dukes that came of the Horites; duke Disan Hus et Aran 36:29 isti duces Horreorum dux Lothan Lotan, duke Shobal, duke Zibeon, duke Anah, 36:30 Duke Dishon, duke Ezer, duke Dishan: these are the dux Sobal dux Sebeon dux Ana 36:30 dux Dison dux Eser dukes that came of Hori, among their dukes in the land of Seir.

dux Disan isti duces Horreorum qui imperauerunt in terra

36:31 And these are the kings that reigned in the land of Seir 36:31 reges autem qui regnauerunt in terra Edom an-

Edom, before there reigned any king over the children of Israel.

tequam haberent regem filii Israhel fuerunt hii 36:32 Bale

36:32 And Bela the son of Beor reigned in Edom: and the name of his city was Dinhabah.

36:33 And Bela died, and Jobab the son of Zerah of filius Beor nomenque urbis eius Denaba 36:33 mortuus est

Bozrah reigned in his stead.

autem Bale et regnauit pro eo Iobab filius Zare de Bosra

36:34 And Jobab died, and Husham of the land of Temani 36:34 cumque mortuus esset Iobab regnauit pro eo Husan de

reigned in his stead.

36:35 And Husham died, and Hadad the son of Bedad, terra Themanorum 36:35 hoc quoque mortuo regnauit pro

who smote Midian in the field of Moab, reigned in his stead: and the name of his city was Avith.

eo Adad filius Badadi qui percussit Madian in regione Moab

36:36 And Hadad died, and Samlah of Masrekah reigned et nomen urbis eius Ahuith 36:36 cumque mortuus esset

in his stead. 36:37 And Samlah died, and Saul of Rehoboth by the river Adad regnauit pro eo Semla de Maserecha 36:37 hoc quoreigned in his stead.

36:38 And Saul died, and Baalhanan the son of Achbor que mortuo regnauit pro eo Saul de fluuio Rooboth 36:38

reigned in his stead.

cumque et hic obisset successit in regnum Baalanam filius

36:39 And Baalhanan the son of Achbor died, and Hadar Achobor 36:39 isto quoque mortuo regnauit pro eo Adad

reigned in his stead: and the name of his city was Pau; and his wife’s name was Mehetabel, the daughter of Matred,

the daughter of Mezahab. nomenque urbis eius Phau et appellabatur uxor illius Me-36:40 And these are the names of the dukes that came ezabel filia Matred filiae Mizaab 36:40 haec ergo nomina

of Esau, according to their families, after their places, by their names; duke Timnah, duke Alvah, duke Jetheth,

Esau in cognationibus et locis et uocabulis suis dux Thamna

36:41 Duke Aholibamah, duke Elah, duke Pinon, dux Alua dux Ietheth 36:41 dux Oolibama dux Ela dux Phi-36:42 Duke Kenaz, duke Teman, duke Mibzar, non 36:42 dux Cenez dux Theman dux Mabsar 36:43 dux

36:43 Duke Magdiel, duke Iram: these be the dukes of Edom, according to their habitations in the land of their possession: he is Esau the father of the Edomites.

Mabdiel dux Iram hii duces Edom habitantes in terra imperii sui ipse est Esau pater Idumeorum

37:1 And Jacob dwelt in the land wherein his father was a 37:1 habitauit autem Iacob in terra Chanaan in qua pereg-

stranger, in the land of Canaan.

37:2 These are the generations of Jacob. Joseph, be- rinatus est pater suus 37:2 et hae sunt generationes eius Io-

ing seventeen years old, was feeding the flock with his brethren; and the lad was with the sons of Bilhah, and with the sons of Zilpah, his father’s wives: and Joseph brought unto his father their evil report.

seph cum sedecim esset annorum pascebat gregem cum fratribus suis adhuc puer et erat cum filiis Balae et Zelphae uxorum patris sui accusauitque fratres suos apud patrem crimine

37:3 Now Israel loved Joseph more than all his children, pessimo 37:3 Israhel autem diligebat Ioseph super omnes

because he was the son of his old age: and he made him a coat of many colours.

Liber Genesis

filios suos eo quod in senectute genuisset eum fecitque ei tu-

nicam polymitam 37:4 uidentes autem fratres eius quod a 37:4 And when his brethren saw that their father loved him more than all his brethren, they hated him, and could not speak peaceably unto him.

patre plus cunctis filiis amaretur oderant eum nec poterant ei quicquam pacificum loqui 37:5 accidit quoque ut uisum 37:5 And Joseph dreamed a dream, and he told it his

brethren: and they hated him yet the more.

somnium referret fratribus quae causa maioris odii seminarium fuit 37:6 dixitque ad eos audite somnium meum quod 37:6 And he said unto them, Hear, I pray you, this dream

which I have dreamed:

uidi 37:7 putabam ligare nos manipulos in agro et quasi 37:7 For, behold, we were binding sheaves in the field,

and, lo, my sheaf arose, and also stood upright; and, behold, your sheaves stood round about, and made obeisance

consurgere manipulum meum et stare uestrosque manipu- to my sheaf.

los circumstantes adorare manipulum meum 37:8 respon- 37:8 And his brethren said to him, Shalt thou indeed reign over us? or shalt thou indeed have dominion over us? And they hated him yet the more for his dreams, and for his

derunt fratres eius numquid rex noster eris aut subiciemur words. dicioni tuae haec ergo causa somniorum atque sermonum inuidiae et odii fomitem ministrauit 37:9 aliud quoque uidit 37:9 And he dreamed yet another dream, and told it his

brethren, and said, Behold, I have dreamed a dream more; and, behold, the sun and the moon and the eleven stars

somnium quod narrans fratribus ait uidi per somnium quasi made obeisance to me.

solem et lunam et stellas undecim adorare me 37:10 quod 37:10 And he told it to his father, and to his brethren: and his father rebuked him, and said unto him, What is this dream that thou hast dreamed? Shall I and thy mother and thy brethren indeed come to bow down ourselves to thee to the earth?

cum patri suo et fratribus rettulisset increpauit eum pater et dixit quid sibi uult hoc somnium quod uidisti num ego et mater tua et fratres adorabimus te super terram 37:11 inui- 37:11 And his brethren envied him; but his father observed

the saying.

debant igitur ei fratres sui pater uero rem tacitus considerabat 37:12 cumque fratres illius in pascendis gregibus patris 37:12 And his brethren went to feed their father’s flock in

Shechem.

morarentur in Sychem 37:13 dixit ad eum Israhel fratres tui 37:13 And Israel said unto Joseph, Do not thy brethren

feed the flock in Shechem? come, and I will send thee unto them. And he said to him, Here am I.

pascunt oues in Sycimis ueni mittam te ad eos quo respondente 37:14 praesto sum ait uade et uide si cuncta prospera 37:14 And he said to him, Go, I pray thee, see whether it

be well with thy brethren, and well with the flocks; and bring me word again. So he sent him out of the vale of

sint erga fratres tuos et pecora et renuntia mihi quid aga- Hebron, and he came to Shechem. tur missus de ualle Hebron uenit in Sychem 37:15 inuenit- 37:15 And a certain man found him, and, behold, he was

wandering in the field: and the man asked him, saying, What seekest thou?

que eum uir errantem in agro et interrogauit quid quaereret 37:16 at ille respondit fratres meos quaero indica mihi ubi pa- 37:16 And he said, I seek my brethren: tell me, I pray thee,

where they feed their flocks.

scant greges 37:17 dixitque ei uir recesserunt de loco isto 37:17 And the man said, They are departed hence; for I

heard them say, Let us go to Dothan. And Joseph went after his brethren, and found them in Dothan.

audiui autem eos dicentes eamus in Dothain perrexit ergo Ioseph post fratres suos et inuenit eos in Dothain 37:18 qui 37:18 And when they saw him afar off, even before he

came near unto them, they conspired against him to slay him.

cum uidissent eum procul antequam accederet ad eos cogitauerunt illum occidere 37:19 et mutuo loquebantur ecce 37:19 And they said one to another, Behold, this dreamer

cometh.

somniator uenit 37:20 uenite occidamus eum et mittamus in 37:20 Come now therefore, and let us slay him, and cast

him into some pit, and we will say, Some evil beast hath devoured him: and we shall see what will become of his

cisternam ueterem dicemusque fera pessima deuorauit eum dreams.

Biblia Sacra Vulgata

37:21 And Reuben heard it, and he delivered him out of et tunc apparebit quid illi prosint somnia sua 37:21 audiens

their hands; and said, Let us not kill him.

hoc Ruben nitebatur liberare eum de manibus eorum et di-

37:22 And Reuben said unto them, Shed no blood, but cebat 37:22 non interficiamus animam eius nec effundatis

cast him into this pit that is in the wilderness, and lay no hand upon him; that he might rid him out of their hands,

to deliver him to his father again. sanguinem sed proicite eum in cisternam hanc quae est in solitudine manusque uestras seruate innoxias hoc autem dicebat uolens eripere eum de manibus eorum et reddere patri

37:23 And it came to pass, when Joseph was come unto his suo 37:23 confestim igitur ut peruenit ad fratres nudauerunt

brethren, that they stript Joseph out of his coat, his coat of many colours that was on him;

eum tunica talari et polymita 37:24 miseruntque in cister-

37:24 And they took him, and cast him into a pit: and the pit was empty, there was no water in it.

37:25 And they sat down to eat bread: and they lifted up nam quae non habebat aquam 37:25 et sedentes ut comede-

their eyes and looked, and, behold, a company of Ishmeelites came from Gilead with their camels bearing spicery

and balm and myrrh, going to carry it down to Egypt. rent panem uiderunt uiatores Ismahelitas uenire de Galaad

et camelos eorum portare aromata et resinam et stacten in

37:26 And Judah said unto his brethren, What profit is it if Aegyptum 37:26 dixit ergo Iudas fratribus suis quid nobis

we slay our brother, and conceal his blood?

prodeet si occiderimus fratrem nostrum et celauerimus san-

37:27 Come, and let us sell him to the Ishmeelites, and let guinem ipsius 37:27 melius est ut uendatur Ismahelitis et

not our hand be upon him; for he is our brother and our flesh. And his brethren were content.

manus nostrae non polluantur frater enim et caro nostra est

37:28 Then there passed by Midianites merchantmen; and adquieuerunt fratres sermonibus eius 37:28 et praetereun-

they drew and lifted up Joseph out of the pit, and sold Joseph to the Ishmeelites for twenty pieces of silver: and

they brought Joseph into Egypt. tibus Madianitis negotiatoribus extrahentes eum de cisterna

uendiderunt Ismahelitis uiginti argenteis qui duxerunt eum

37:29 And Reuben returned unto the pit; and, behold, in Aegyptum 37:29 reuersusque Ruben ad cisternam non

Joseph was not in the pit; and he rent his clothes.

37:30 And he returned unto his brethren, and said, The inuenit puerum 37:30 et scissis uestibus pergens ad fratres

child is not; and I, whither shall I go? 37:31 And they took Joseph’s coat, and killed a kid of the ait puer non conparet et ego quo ibo 37:31 tulerunt autem goats, and dipped the coat in the blood;

tunicam eius et in sanguinem hedi quem occiderant tinxerunt

37:32 And they sent the coat of many colours, and they 37:32 mittentes qui ferrent ad patrem et dicerent hanc inueni-

brought it to their father; and said, This have we found: know now whether it be thy son’s coat or no.

mus uide utrum tunica filii tui sit an non 37:33 quam cum

37:33 And he knew it, and said, It is my son’s coat; an evil beast hath devoured him; Joseph is without doubt rent in pieces.

agnouisset pater ait tunica filii mei est fera pessima come-

37:34 And Jacob rent his clothes, and put sackcloth upon dit eum bestia deuorauit Ioseph 37:34 scissisque uestibus

his loins, and mourned for his son many days.

37:35 And all his sons and all his daughters rose up to indutus est cilicio lugens filium multo tempore 37:35 con-

comfort him; but he refused to be comforted; and he said, For I will go down into the grave unto my son mourning.

Thus his father wept for him. gregatis autem cunctis liberis eius ut lenirent dolorem pat-

ris noluit consolationem recipere et ait descendam ad filium

37:36 And the Midianites sold him into Egypt unto meum lugens in infernum et illo perseuerante in fletu 37:36

Potiphar, an officer of Pharaoh’s, and captain of the guard.

Madianei uendiderunt Ioseph in Aegypto Putiphar eunucho Pharaonis magistro militiae

38:1 And it came to pass at that time, that Judah went 38:1 eo tempore descendens Iudas a fratribus suis diuer-

down from his brethren, and turned in to a certain Adullamite, whose name was Hirah.

Liber Genesis

tit ad uirum odollamitem nomine Hiram 38:2 uiditque ibi 38:2 And Judah saw there a daughter of a certain Canaan-

ite, whose name was Shuah; and he took her, and went in unto her.

filiam hominis chananei uocabulo Suae et uxore accepta ingressus est ad eam 38:3 quae concepit et peperit filium uo- 38:3 And she conceived, and bare a son; and he called his

name Er.

cauitque nomen eius Her 38:4 rursum concepto fetu natum 38:4 And she conceived again, and bare a son; and she

called his name Onan.

filium nominauit Onam 38:5 tertium quoque peperit quem 38:5 And she yet again conceived, and bare a son; and

called his name Shelah: and he was at Chezib, when she bare him.

appellauit Sela quo nato parere ultra cessauit 38:6 dedit

38:6 And Judah took a wife for Er his firstborn, whose name was Tamar.

autem Iudas uxorem primogenito suo Her nomine Thamar 38:7 fuitque Her primogenitus Iudae nequam in conspectu 38:7 And Er, Judah’s firstborn, was wicked in the sight of

the LORD; and the LORD slew him.

Domini et ab eo occisus est 38:8 dixit ergo Iudas ad Onam 38:8 And Judah said unto Onan, Go in unto thy brother’s

wife, and marry her, and raise up seed to thy brother.

filium suum ingredere ad uxorem fratris tui et sociare illi ut suscites semen fratri tuo 38:9 ille sciens non sibi nasci fi- 38:9 And Onan knew that the seed should not be his; and

it came to pass, when he went in unto his brother’s wife, that he spilled it on the ground, lest that he should give

lios introiens ad uxorem fratris sui semen fundebat in terram seed to his brother. ne liberi fratris nomine nascerentur 38:10 et idcirco per- 38:10 And the thing which he did displeased the LORD:

wherefore he slew him also.

cussit eum Dominus quod rem detestabilem faceret 38:11 38:11 Then said Judah to Tamar his daughter in law, Re-

main a widow at thy father’s house, till Shelah my son be grown: for he said, Lest peradventure he die also, as his brethren did. And Tamar went and dwelt in her father’s house.

quam ob rem dixit Iudas Thamar nurui suae esto uidua in domo patris tui donec crescat Sela filius meus timebat enim ne et ipse moreretur sicut fratres eius quae abiit et habitauit in domo patris sui 38:12 euolutis autem multis diebus 38:12 And in process of time the daughter of Shuah Ju-

dah’s wife died; and Judah was comforted, and went up unto his sheepshearers to Timnath, he and his friend Hirah

mortua est filia Suae uxor Iudae qui post luctum consola- the Adullamite. tione suscepta ascendebat ad tonsores ouium suarum ipse et Hiras opilio gregis Odollamita in Thamnas 38:13 nuntia- 38:13 And it was told Tamar, saying, Behold thy father in

law goeth up to Timnath to shear his sheep.

tumque est Thamar quod socer illius ascenderet in Thamnas ad tondendas oues 38:14 quae depositis uiduitatis uestibus 38:14 And she put her widow’s garments off from her, and

covered her with a vail, and wrapped herself, and sat in an open place, which is by the way to Timnath; for she saw that Shelah was grown, and she was not given unto him to wife.

adsumpsit theristrum et mutato habitu sedit in biuio itineris quod ducit Thamnam eo quod creuisset Sela et non eum accepisset maritum 38:15 quam cum uidisset Iudas suspicatus 38:15 When Judah saw her, he thought her to be an harlot;

because she had covered her face.

est esse meretricem operuerat enim uultum suum ne cognosceretur 38:16 ingrediensque ad eam ait dimitte me ut coeam 38:16 And he turned unto her by the way, and said, Go to,

I pray thee, let me come in unto thee; (for he knew not that she was his daughter in law.) And she said, What wilt thou

tecum nesciebat enim quod nurus sua esset qua respondente give me, that thou mayest come in unto me? quid mihi dabis ut fruaris concubitu meo 38:17 dixit mit- 38:17 And he said, I will send thee a kid from the flock.

And she said, Wilt thou give me a pledge, till thou send it?

tam tibi hedum de gregibus rursum illa dicente patiar quod uis si dederis mihi arrabonem donec mittas quod polliceris 38:18 ait Iudas quid uis tibi pro arrabone dari respondit anu- 38:18 And he said, What pledge shall I give thee? And she

said, Thy signet, and thy bracelets, and thy staff that is in thine hand. And he gave it her, and came in unto her, and she conceived by him.

Biblia Sacra Vulgata

lum tuum et armillam et baculum quem manu tenes ad unum

38:19 And she arose, and went away, and laid by her vail igitur coitum concepit mulier 38:19 et surgens abiit deposi-

from her, and put on the garments of her widowhood.

toque habitu quem adsumpserat induta est uiduitatis uestibus

38:20 And Judah sent the kid by the hand of his friend the 38:20 misit autem Iudas hedum per pastorem suum Odolla-

Adullamite, to receive his pledge from the woman’s hand: but he found her not.

mitem ut reciperet pignus quod dederat mulieri qui cum non

38:21 Then he asked the men of that place, saying, Where inuenisset eam 38:21 interrogauit homines loci illius ubi est

is the harlot, that was openly by the way side? And they said, There was no harlot in this place.

mulier quae sedebat in biuio respondentibus cunctis non fuit

38:22 And he returned to Judah, and said, I cannot find in loco isto meretrix 38:22 reuersus est ad Iudam et dixit

her; and also the men of the place said, that there was no harlot in this place.

ei non inueni eam sed et homines loci illius dixerunt mihi

38:23 And Judah said, Let her take it to her, lest we be numquam ibi sedisse scortum 38:23 ait Iudas habeat sibi

shamed: behold, I sent this kid, and thou hast not found her.

certe mendacii nos arguere non poterit ego misi hedum quem

38:24 And it came to pass about three months after, that promiseram et tu non inuenisti eam 38:24 ecce autem post

it was told Judah, saying, Tamar thy daughter in law hath played the harlot; and also, behold, she is with child by whoredom. And Judah said, Bring her forth, and let her be burnt.

tres menses nuntiauerunt Iudae dicentes fornicata est Thamar nurus tua et uidetur uterus illius intumescere dixit Iudas

38:25 When she was brought forth, she sent to her father producite eam ut conburatur 38:25 quae cum educeretur ad

in law, saying, By the man, whose these are, am I with child: and she said, Discern, I pray thee, whose are these,

the signet, and bracelets, and staff. poenam misit ad socerum suum dicens de uiro cuius haec

sunt concepi cognosce cuius sit anulus et armilla et bacu-

38:26 And Judah acknowledged them, and said, She hath lus 38:26 qui agnitis muneribus ait iustior me est quia non

been more righteous than I; because that I gave her not to Shelah my son. And he knew her again no more.

tradidi eam Sela filio meo attamen ultra non cognouit illam

38:27 And it came to pass in the time of her travail, that, 38:27 instante autem partu apparuerunt gemini in utero at-

behold, twins were in her womb.

que in ipsa effusione infantum unus protulit manum in qua

38:28 And it came to pass, when she travailed, that the one obsetrix ligauit coccinum dicens 38:28 iste egreditur prior

put out his hand: and the midwife took and bound upon his hand a scarlet thread, saying, This came out first.

38:29 illo uero retrahente manum egressus est alter dixitque

38:29 And it came to pass, as he drew back his hand, that, behold, his brother came out: and she said, How hast thou broken forth? this breach be upon thee: therefore his name was called Pharez.

mulier quare diuisa est propter te maceria et ob hanc causam

38:30 And afterward came out his brother, that had the uocauit nomen eius Phares 38:30 postea egressus est frater

scarlet thread upon his hand: and his name was called Zarah.

in cuius manu erat coccinum quem appellauit Zara

39:1 And Joseph was brought down to Egypt; and 39:1 igitur Ioseph ductus est in Aegyptum emitque eum

Potiphar, an officer of Pharaoh, captain of the guard, an Egyptian, bought him of the hands of the Ishmeelites,

which had brought him down thither. Putiphar eunuchus Pharaonis princeps exercitus uir aegyp-39:2 And the LORD was with Joseph, and he was a pros- tius de manu Ismahelitarum a quibus perductus erat 39:2

perous man; and he was in the house of his master the Egyptian.

fuitque Dominus cum eo et erat uir in cunctis prospere agens

39:3 And his master saw that the LORD was with him, habitabatque in domo domini sui 39:3 qui optime nouerat

and that the LORD made all that he did to prosper in his hand.

esse Dominum cum eo et omnia quae gereret ab eo dirigi

39:4 And Joseph found grace in his sight, and he served in manu illius 39:4 inuenitque Ioseph gratiam coram do-

him: and he made him overseer over his house, and all that he had he put into his hand.

Liber Genesis

mino suo et ministrabat ei a quo praepositus omnibus gubernabat creditam sibi domum et uniuersa quae tradita fuerant 39:5 benedixitque Dominus domui Aegyptii propter Ioseph 39:5 And it came to pass from the time that he had made

him overseer in his house, and over all that he had, that the LORD blessed the Egyptian’s house for Joseph’s sake; and the blessing of the LORD was upon all that he had in the house, and in the field.

et multiplicauit tam in aedibus quam in agris cunctam eius substantiam 39:6 nec quicquam aliud nouerat nisi panem 39:6 And he left all that he had in Joseph’s hand; and he

knew not ought he had, save the bread which he did eat. And Joseph was a goodly person, and well favoured.

quo uescebatur erat autem Ioseph pulchra facie et decorus aspectu 39:7 post multos itaque dies iecit domina oculos 39:7 And it came to pass after these things, that his mas-

ter’s wife cast her eyes upon Joseph; and she said, Lie with me.

suos in Ioseph et ait dormi mecum 39:8 qui nequaquam ad-

39:8 But he refused, and said unto his master’s wife, Behold, my master wotteth not what is with me in the house, and he hath committed all that he hath to my hand;

quiescens operi nefario dixit ad eam ecce dominus meus omnibus mihi traditis ignorat quid habeat in domo sua 39:9 nec 39:9 There is none greater in this house than I; neither hath

he kept back any thing from me but thee, because thou art his wife: how then can I do this great wickedness, and sin

quicquam est quod non in mea sit potestate uel non tradide- against God? rit mihi praeter te quae uxor eius es quomodo ergo possum malum hoc facere et peccare in Deum meum 39:10 huius- 39:10 And it came to pass, as she spake to Joseph day by

day, that he hearkened not unto her, to lie by her, or to be with her.

cemodi uerbis per singulos dies et mulier molesta erat adulescenti et ille recusabat stuprum 39:11 accidit autem ut qua- 39:11 And it came to pass about this time, that Joseph

went into the house to do his business; and there was none of the men of the house there within.

dam die intraret Ioseph domum et operis quippiam absque arbitris faceret 39:12 et illa adprehensa lacinia uestimenti 39:12 And she caught him by his garment, saying, Lie

with me: and he left his garment in her hand, and fled, and got him out.

eius diceret dormi mecum qui relicto in manu illius pallio fugit et egressus est foras 39:13 cumque uidisset mulier ue- 39:13 And it came to pass, when she saw that he had left

his garment in her hand, and was fled forth,

stem in manibus suis et se esse contemptam 39:14 uocauit 39:14 That she called unto the men of her house, and spake

unto them, saying, See, he hath brought in an Hebrew unto us to mock us; he came in unto me to lie with me, and I

homines domus suae et ait ad eos en introduxit uirum heb- cried with a loud voice: raeum ut inluderet nobis ingressus est ad me ut coiret mecum cumque ego succlamassem 39:15 et audisset uocem meam 39:15 And it came to pass, when he heard that I lifted up

my voice and cried, that he left his garment with me, and fled, and got him out.

reliquit pallium quod tenebam et fugit foras 39:16 in argu-

39:16 And she laid up his garment by her, until his lord came home.

mentum ergo fidei retentum pallium ostendit marito reuertenti domum 39:17 et ait ingressus est ad me seruus heb- 39:17 And she spake unto him according to these words,

saying, The Hebrew servant, which thou hast brought unto us, came in unto me to mock me:

raeus quem adduxisti ut inluderet mihi 39:18 cumque uidis-

39:18 And it came to pass, as I lifted up my voice and cried, that he left his garment with me, and fled out.

set me clamare reliquit pallium et fugit foras 39:19 his au- _{39:19 And it came to pass, when his master heard the}

words of his wife, which she spake unto him, saying, After this manner did thy servant to me; that his wrath was

ditis dominus et nimium credulus uerbis coniugis iratus est _{kindled.} ualde 39:20 tradiditque Ioseph in carcerem ubi uincti regis 39:20 And Joseph’s master took him, and put him into the

prison, a place where the king’s prisoners were bound: and he was there in the prison.

custodiebantur et erat ibi clausus 39:21 fuit autem Dominus

39:21 But the LORD was with Joseph, and shewed him mercy, and gave him favour in the sight of the keeper of the prison.

cum Ioseph et misertus illius dedit ei gratiam in conspectu principis carceris 39:22 qui tradidit in manu ipsius uniuer- 39:22 And the keeper of the prison committed to Joseph’s

hand all the prisoners that were in the prison; and whatsoever they did there, he was the doer of it.

Biblia Sacra Vulgata

sos uinctos qui in custodia tenebantur et quicquid fiebat sub

39:23 The keeper of the prison looked not to any thing that ipso erat 39:23 nec nouerat aliquid cunctis ei creditis Do-

was under his hand; because the LORD was with him, and that which he did, the LORD made it to prosper.

minus enim erat cum illo et omnia eius opera dirigebat

40:1 And it came to pass after these things, that the butler 40:1 his ita gestis accidit ut peccarent duo eunuchi pinof the king of Egypt and his baker had offended their lord the king of Egypt.

cerna regis Aegypti et pistor domino suo 40:2 iratusque

40:2 And Pharaoh was wroth against two of his officers, against the chief of the butlers, and against the chief of the bakers.

Pharao contra eos nam alter pincernis praeerat alter pisto-

40:3 And he put them in ward in the house of the captain ribus 40:3 misit eos in carcerem principis militum in quo

of the guard, into the prison, the place where Joseph was bound.

erat uinctus et Ioseph 40:4 at custos carceris tradidit eos

40:4 And the captain of the guard charged Joseph with them, and he served them: and they continued a season in ward.

Ioseph qui et ministrabat eis aliquantum temporis fluxerat

40:5 And they dreamed a dream both of them, each man et illi in custodia tenebantur 40:5 uideruntque ambo som-

his dream in one night, each man according to the interpretation of his dream, the butler and the baker of the king

of Egypt, which were bound in the prison. nium nocte una iuxta interpretationem congruam sibi 40:6

40:6 And Joseph came in unto them in the morning, and looked upon them, and, behold, they were sad.

ad quos cum introisset Ioseph mane et uidisset eos tristes

40:7 And he asked Pharaoh’s officers that were with him 40:7 sciscitatus est dicens cur tristior est hodie solito facies

in the ward of his lord’s house, saying, Wherefore look ye so sadly to day?

uestra 40:8 qui responderunt somnium uidimus et non est

40:8 And they said unto him, We have dreamed a dream, and there is no interpreter of it. And Joseph said unto them, Do not interpretations belong to God? tell me them, I pray you.

qui interpretetur nobis dixitque ad eos Ioseph numquid non

40:9 And the chief butler told his dream to Joseph, and Dei est interpretatio referte mihi quid uideritis 40:9 narrauit

said to him, In my dream, behold, a vine was before me;

prior praepositus pincernarum somnium uidebam coram me

40:10 And in the vine were three branches: and it was uitem 40:10 in qua erant tres propagines crescere paulatim

as though it budded, and her blossoms shot forth; and the clusters thereof brought forth ripe grapes:

gemmas et post flores uuas maturescere 40:11 calicemque

40:11 And Pharaoh’s cup was in my hand: and I took the grapes, and pressed them into Pharaoh’s cup, and I gave the cup into Pharaoh’s hand.

Pharaonis in manu mea tuli ergo uuas et expressi in calicem

40:12 And Joseph said unto him, This is the interpretation quem tenebam et tradidi poculum Pharaoni 40:12 responof it: The three branches are three days:

dit Ioseph haec est interpretatio somnii tres propagines tres

40:13 Yet within three days shall Pharaoh lift up thine adhuc dies sunt 40:13 post quos recordabitur Pharao ma-

head, and restore thee unto thy place: and thou shalt deliver Pharaoh’s cup into his hand, after the former manner

when thou wast his butler. gisterii tui et restituet te in gradum pristinum dabisque ei

calicem iuxta officium tuum sicut facere ante consueueras

40:14 But think on me when it shall be well with thee, and 40:14 tantum memento mei cum tibi bene fuerit et facies me-

shew kindness, I pray thee, unto me, and make mention of me unto Pharaoh, and bring me out of this house:

cum misericordiam ut suggeras Pharaoni et educat me de

40:15 For indeed I was stolen away out of the land of isto carcere 40:15 quia furto sublatus sum de terra Hebrae-

the Hebrews: and here also have I done nothing that they should put me into the dungeon.

orum et hic innocens in lacum missus sum 40:16 uidens pi-

40:16 When the chief baker saw that the interpretation was good, he said unto Joseph, I also was in my dream, and, behold, I had three white baskets on my head:

storum magister quod prudenter somnium dissoluisset ait et ego uidi somnium quod haberem tria canistra farinae super

40:17 And in the uppermost basket there was of all manner caput meum 40:17 et in uno canistro quod erat excelsius of bakemeats for Pharaoh; and the birds did eat them out of the basket upon my head.

portare me omnes cibos qui fiunt arte pistoria auesque co-

Liber Genesis

medere ex eo 40:18 respondit Ioseph haec est interpretatio 40:18 And Joseph answered and said, This is the interpre-

tation thereof: The three baskets are three days:

somnii tria canistra tres adhuc dies sunt 40:19 post quos au- 40:19 Yet within three days shall Pharaoh lift up thy head

from off thee, and shall hang thee on a tree; and the birds shall eat thy flesh from off thee.

feret Pharao caput tuum ac suspendet te in cruce et lacerabunt uolucres carnes tuas 40:20 exin dies tertius natalicius 40:20 And it came to pass the third day, which was

Pharaoh’s birthday, that he made a feast unto all his servants: and he lifted up the head of the chief butler and of

Pharaonis erat qui faciens grande conuiuium pueris suis re- the chief baker among his servants. cordatus est inter epulas magistri pincernarum et pistorum principis 40:21 restituitque alterum in locum suum ut por- 40:21 And he restored the chief butler unto his butlership

again; and he gave the cup into Pharaoh’s hand:

rigeret regi poculum 40:22 alterum suspendit in patibulo ut 40:22 But he hanged the chief baker: as Joseph had inter-

preted to them.

coniectoris ueritas probaretur 40:23 et tamen succedentibus 40:23 Yet did not the chief butler remember Joseph, but

forgat him.

prosperis praepositus pincernarum oblitus est interpretis sui

41:1 post duos annos uidit Pharao somnium putabat se 41:1 And it came to pass at the end of two full years, that

Pharaoh dreamed: and, behold, he stood by the river.

stare super fluuium 41:2 de quo ascendebant septem boues 41:2 And, behold, there came up out of the river seven well

favoured kine and fatfleshed; and they fed in a meadow.

pulchrae et crassae nimis et pascebantur in locis palustribus 41:3 aliae quoque septem emergebant de flumine foedae con- 41:3 And, behold, seven other kine came up after them out

of the river, ill favoured and leanfleshed; and stood by the other kine upon the brink of the river.

fectaeque macie et pascebantur in ipsa amnis ripa in locis uirentibus 41:4 deuoraueruntque eas quarum mira species 41:4 And the ill favoured and leanfleshed kine did eat up

the seven well favoured and fat kine. So Pharaoh awoke.

et habitudo corporum erat expergefactus Pharao 41:5 rur- 41:5 And he slept and dreamed the second time: and, be-

hold, seven ears of corn came up upon one stalk, rank and good.

sum dormiuit et uidit alterum somnium septem spicae pul-

lulabant in culmo uno plenae atque formonsae 41:6 aliae 41:6 And, behold, seven thin ears and blasted with the east

wind sprung up after them.

quoque totidem spicae tenues et percussae uredine oriebantur 41:7 deuorantes omnem priorum pulchritudinem euigi- 41:7 And the seven thin ears devoured the seven rank and

full ears. And Pharaoh awoke, and, behold, it was a dream.

lans post quietem 41:8 et facto mane pauore perterritus mi- 41:8 And it came to pass in the morning that his spirit

was troubled; and he sent and called for all the magicians of Egypt, and all the wise men thereof: and Pharaoh told them his dream; but there was none that could interpret them unto Pharaoh.

sit ad coniectores Aegypti cunctosque sapientes et accersitis narrauit somnium nec erat qui interpretaretur 41:9 tunc 41:9 Then spake the chief butler unto Pharaoh, saying, I

do remember my faults this day:

demum reminiscens pincernarum magister ait confiteor peccatum meum 41:10 iratus rex seruis suis me et magistrum 41:10 Pharaoh was wroth with his servants, and put me in

ward in the captain of the guard’s house, both me and the chief baker:

pistorum retrudi iussit in carcerem principis militum 41:11

41:11 And we dreamed a dream in one night, I and he; we dreamed each man according to the interpretation of his dream.

ubi una nocte uterque uidimus somnium praesagum futurorum 41:12 erat ibi puer hebraeus eiusdem ducis militum fa- 41:12 And there was there with us a young man, an He-

brew, servant to the captain of the guard; and we told him, and he interpreted to us our dreams; to each man accord-

mulus cui narrantes somnia 41:13 audiuimus quicquid post- ing to his dream he did interpret.

41:13 And it came to pass, as he interpreted to us, so it was; me he restored unto mine office, and him he hanged.

ea rei probauit euentus ego enim redditus sum officio meo et ille suspensus est in cruce 41:14 protinus ad regis imperium 41:14 Then Pharaoh sent and called Joseph, and they

brought him hastily out of the dungeon: and he shaved himself, and changed his raiment, and came in unto

eductum de carcere Ioseph totonderunt ac ueste mutata obtu- Pharaoh.

Biblia Sacra Vulgata

41:15 And Pharaoh said unto Joseph, I have dreamed a lerunt ei 41:15 cui ille ait uidi somnia nec est qui edisserat

dream, and there is none that can interpret it: and I have heard say of thee, that thou canst understand a dream to

interpret it. quae audiui te prudentissime conicere 41:16 respondit Io-

41:16 And Joseph answered Pharaoh, saying, It is not in me: God shall give Pharaoh an answer of peace.

seph absque me Deus respondebit prospera Pharaoni 41:17

41:17 And Pharaoh said unto Joseph, In my dream, behold, I stood upon the bank of the river:

narrauit ergo ille quod uiderat putabam me stare super ri-

41:18 And, behold, there came up out of the river seven pam fluminis 41:18 et septem boues de amne conscendere kine, fatfleshed and well favoured; and they fed in a meadow:

pulchras nimis et obesis carnibus quae in pastu paludis ui-

41:19 And, behold, seven other kine came up after them, recta carpebant 41:19 et ecce has sequebantur aliae septem

poor and very ill favoured and leanfleshed, such as I never saw in all the land of Egypt for badness:

boues in tantum deformes et macilentae ut numquam tales

41:20 And the lean and the ill favoured kine did eat up the in terra Aegypti uiderim 41:20 quae deuoratis et consump-

first seven fat kine:

41:21 And when they had eaten them up, it could not be tis prioribus 41:21 nullum saturitatis dedere uestigium sed

known that they had eaten them; but they were still ill favoured, as at the beginning. So I awoke.

simili macie et squalore torpebant euigilans rursum sopore

41:22 And I saw in my dream, and, behold, seven ears depressus 41:22 uidi somnium septem spicae pullulabant

came up in one stalk, full and good: 41:23 And, behold, seven ears, withered, thin, and blasted in culmo uno plenae atque pulcherrimae 41:23 aliae quowith the east wind, sprung up after them:

que septem tenues et percussae uredine oriebantur stipula

41:24 And the thin ears devoured the seven good ears: 41:24 quae priorum pulchritudinem deuorarunt narraui con-

and I told this unto the magicians; but there was none that could declare it to me.

iectoribus somnium et nemo est qui edisserat 41:25 respon-

41:25 And Joseph said unto Pharaoh, The dream of Pharaoh is one: God hath shewed Pharaoh what he is about to do.

dit Ioseph somnium regis unum est quae facturus est Deus

41:26 The seven good kine are seven years; and the seven ostendit Pharaoni 41:26 septem boues pulchrae et septem

good ears are seven years: the dream is one.

spicae plenae septem ubertatis anni sunt eandemque uim

41:27 And the seven thin and ill favoured kine that came somnii conprehendunt 41:27 septem quoque boues tenues

up after them are seven years; and the seven empty ears blasted with the east wind shall be seven years of famine.

atque macilentae quae ascenderunt post eas et septem spicae tenues et uento urente percussae septem anni sunt uen-

41:28 This is the thing which I have spoken unto Pharaoh: turae famis 41:28 qui hoc ordine conplebuntur 41:29 ecce

What God is about to do he sheweth unto Pharaoh. 41:29 Behold, there come seven years of great plenty throughout all the land of Egypt:

septem anni uenient fertilitatis magnae in uniuersa terra Ae-

41:30 And there shall arise after them seven years of gypti 41:30 quos sequentur septem anni alii tantae sterili-

famine; and all the plenty shall be forgotten in the land of Egypt; and the famine shall consume the land;

tatis ut obliuioni tradatur cuncta retro abundantia consump-

41:31 And the plenty shall not be known in the land by rea- tura est enim fames omnem terram 41:31 et ubertatis mason of that famine following; for it shall be very grievous.

41:32 And for that the dream was doubled unto Pharaoh gnitudinem perditura inopiae magnitudo 41:32 quod autem

twice; it is because the thing is established by God, and God will shortly bring it to pass.

uidisti secundo ad eandem rem pertinens somnium firmitatis indicium est eo quod fiat sermo Dei et uelocius implea-

41:33 Now therefore let Pharaoh look out a man discreet tur 41:33 nunc ergo prouideat rex uirum sapientem et indu-

and wise, and set him over the land of Egypt.

41:34 Let Pharaoh do this, and let him appoint officers strium et praeficiat eum terrae Aegypti 41:34 qui constituat

over the land, and take up the fifth part of the land of Egypt in the seven plenteous years.

praepositos per singulas regiones et quintam partem fruc-

Liber Genesis

tuum per septem annos fertilitatis 41:35 qui iam nunc fu- 41:35 And let them gather all the food of those good years

that come, and lay up corn under the hand of Pharaoh, and let them keep food in the cities.

turi sunt congreget in horrea et omne frumentum sub Pharaonis potestate condatur serueturque in urbibus 41:36 et 41:36 And that food shall be for store to the land against

the seven years of famine, which shall be in the land of Egypt; that the land perish not through the famine.

paretur futurae septem annorum fami quae pressura est Aegyptum et non consumetur terra inopia 41:37 placuit Pha- 41:37 And the thing was good in the eyes of Pharaoh, and

in the eyes of all his servants.

raoni consilium et cunctis ministris eius 41:38 locutusque 41:38 And Pharaoh said unto his servants, Can we find

such a one as this is, a man in whom the Spirit of God is?

est ad eos num inuenire poterimus talem uirum qui spiritu Dei plenus sit 41:39 dixit ergo ad Ioseph quia ostendit De- 41:39 And Pharaoh said unto Joseph, Forasmuch as God

hath shewed thee all this, there is none so discreet and wise as thou art:

us tibi omnia quae locutus es numquid sapientiorem et similem tui inuenire potero 41:40 tu eris super domum meam 41:40 Thou shalt be over my house, and according unto

thy word shall all my people be ruled: only in the throne will I be greater than thou.

et ad tui oris imperium cunctus populus oboediet uno tantum regni solio te praecedam 41:41 dicens quoque rursum 41:41 And Pharaoh said unto Joseph, See, I have set thee

over all the land of Egypt.

Pharao ad Ioseph ecce constitui te super uniuersam terram Aegypti 41:42 tulit anulum de manu sua et dedit in manu 41:42 And Pharaoh took off his ring from his hand, and

put it upon Joseph’s hand, and arrayed him in vestures of fine linen, and put a gold chain about his neck;

eius uestiuitque eum stola byssina et collo torquem auream circumposuit 41:43 fecitque ascendere super currum suum 41:43 And he made him to ride in the second chariot which

he had; and they cried before him, Bow the knee: and he made him ruler over all the land of Egypt.

secundum clamante praecone ut omnes coram eo genuflecterent et praepositum esse scirent uniuersae terrae Aegypti 41:44 dixit quoque rex ad Ioseph ego sum Pharao absque tuo 41:44 And Pharaoh said unto Joseph, I am Pharaoh, and

without thee shall no man lift up his hand or foot in all the land of Egypt.

imperio non mouebit quisquam manum aut pedem in omni terra Aegypti 41:45 uertitque nomen illius et uocauit eum 41:45 And Pharaoh called Joseph’s name Zaphnath-

paaneah; and he gave him to wife Asenath the daughter of Potipherah priest of On. And Joseph went out over all

lingua aegyptiaca Saluatorem mundi dedit quoque illi uxo- the land of Egypt. rem Aseneth filiam Putiphare sacerdotis Heliopoleos egressus itaque Ioseph ad terram Aegypti 41:46 triginta autem 41:46 And Joseph was thirty years old when he stood be-

fore Pharaoh king of Egypt. And Joseph went out from the presence of Pharaoh, and went throughout all the land

erat annorum quando stetit in conspectu regis Pharaonis cir- of Egypt. cuiuit omnes regiones Aegypti 41:47 uenitque fertilitas se- 41:47 And in the seven plenteous years the earth brought

forth by handfuls.

ptem annorum et in manipulos redactae segetes congregatae sunt in horrea Aegypti 41:48 omnis etiam frugum abundan- 41:48 And he gathered up all the food of the seven years,

which were in the land of Egypt, and laid up the food in the cities: the food of the field, which was round about

tia in singulis urbibus condita est 41:49 tantaque fuit mul- every city, laid he up in the same.

41:49 And Joseph gathered corn as the sand of the sea, very much, until he left numbering; for it was without number.

titudo tritici ut harenae maris coaequaretur et copia mensuram excederet 41:50 nati sunt autem Ioseph filii duo ante- 41:50 And unto Joseph were born two sons before the

years of famine came, which Asenath the daughter of Potipherah priest of On bare unto him.

quam ueniret fames quos ei peperit Aseneth filia Putiphare sacerdotis Heliopoleos 41:51 uocauitque nomen primoge- 41:51 And Joseph called the name of the firstborn Man-

asseh: For God, said he, hath made me forget all my toil, and all my father’s house.

Biblia Sacra Vulgata

niti Manasse dicens obliuisci me fecit Deus omnium labo-

41:52 And the name of the second called he Ephraim: For rum meorum et domum patris mei 41:52 nomen quoque God hath caused me to be fruitful in the land of my affliction.

secundi appellauit Ephraim dicens crescere me fecit Deus

41:53 And the seven years of plenteousness, that was in in terra paupertatis meae 41:53 igitur transactis septem an-

the land of Egypt, were ended.

41:54 And the seven years of dearth began to come, ac- nis ubertatis qui fuerant in Aegypto 41:54 coeperunt uenire

cording as Joseph had said: and the dearth was in all lands; but in all the land of Egypt there was bread.

septem anni inopiae quos praedixerat Ioseph et in uniuerso orbe fames praeualuit in cuncta autem terra Aegypti erat pa-

41:55 And when all the land of Egypt was famished, the nis 41:55 qua esuriente clamauit populus ad Pharaonem ali-

people cried to Pharaoh for bread: and Pharaoh said unto all the Egyptians, Go unto Joseph; what he saith to you,

do. menta petens quibus ille respondit ite ad Ioseph et quicquid 41:56 And the famine was over all the face of the earth: uobis dixerit facite 41:56 crescebat autem cotidie fames in

and Joseph opened all the storehouses, and sold unto the Egyptians; and the famine waxed sore in the land of Egypt.

omni terra aperuitque Ioseph uniuersa horrea et uendebat

41:57 And all countries came into Egypt to Joseph for to Aegyptiis nam et illos oppresserat fames 41:57 omnesque

buy corn; because that the famine was so sore in all lands.

prouinciae ueniebant in Aegyptum ut emerent escas et malum inopiae temperarent

42:1 Now when Jacob saw that there was corn in Egypt, 42:1 audiens autem Iacob quod alimenta uenderentur in

Jacob said unto his sons, Why do ye look one upon another?

Aegypto dixit filiis suis quare neglegitis 42:2 audiui quod

42:2 And he said, Behold, I have heard that there is corn in Egypt: get you down thither, and buy for us from thence; that we may live, and not die.

triticum uenundetur in Aegypto descendite et emite nobis necessaria ut possimus uiuere et non consumamur inopia

42:3 And Joseph’s ten brethren went down to buy corn in 42:3 descendentes igitur fratres Ioseph decem ut emerent fru-

Egypt. 42:4 But Benjamin, Joseph’s brother, Jacob sent not with menta in Aegypto 42:4 Beniamin domi retento ab Iacob qui his brethren; for he said, Lest peradventure mischief befall him.

dixerat fratribus eius ne forte in itinere quicquam patiatur

42:5 And the sons of Israel came to buy corn among those mali 42:5 ingressi sunt terram Aegypti cum aliis qui perge-

that came: for the famine was in the land of Canaan.

42:6 And Joseph was the governor over the land, and he bant ad emendum erat autem fames in terra Chanaan 42:6 et

it was that sold to all the people of the land: and Joseph’s brethren came, and bowed down themselves before him

with their faces to the earth. Ioseph princeps Aegypti atque ad illius nutum frumenta po-42:7 And Joseph saw his brethren, and he knew them, but pulis uendebantur cumque adorassent eum fratres sui 42:7

made himself strange unto them, and spake roughly unto them; and he said unto them, Whence come ye? And they

said, From the land of Canaan to buy food. et agnouisset eos quasi ad alienos durius loquebatur inter-

rogans eos unde uenistis qui responderunt de terra Chanaan

42:8 And Joseph knew his brethren, but they knew not ut emamus uictui necessaria 42:8 et tamen fratres ipse cog-

him.

42:9 And Joseph remembered the dreams which he noscens non est agnitus ab eis 42:9 recordatusque somnio-

dreamed of them, and said unto them, Ye are spies; to see the nakedness of the land ye are come.

rum quae aliquando uiderat ait exploratores estis ut uideatis

42:10 And they said unto him, Nay, my lord, but to buy infirmiora terrae uenistis 42:10 qui dixerunt non est ita do-

food are thy servants come.

42:11 We are all one man’s sons; we are true men, thy mine sed serui tui uenerunt ut emerent cibos 42:11 omnes

servants are no spies.

filii unius uiri sumus pacifici uenimus nec quicquam famuli

Liber Genesis

tui machinantur mali 42:12 quibus ille respondit aliter est 42:12 And he said unto them, Nay, but to see the naked-

ness of the land ye are come.

inmunita terrae huius considerare uenistis 42:13 et illi duo- 42:13 And they said, Thy servants are twelve brethren, the

sons of one man in the land of Canaan; and, behold, the youngest is this day with our father, and one is not.

decim inquiunt serui tui fratres sumus filii uiri unius in terra Chanaan minimus cum patre nostro est alius non est super 42:14 hoc est ait quod locutus sum exploratores estis 42:15 42:14 And Joseph said unto them, That is it that I spake

unto you, saying, Ye are spies:

42:15 Hereby ye shall be proved: By the life of Pharaoh ye shall not go forth hence, except your youngest brother come hither.

iam nunc experimentum uestri capiam per salutem Pharaonis non egrediemini hinc donec ueniat frater uester minimus 42:16 mittite e uobis unum et adducat eum uos autem eritis 42:16 Send one of you, and let him fetch your brother, and

ye shall be kept in prison, that your words may be proved, whether there be any truth in you: or else by the life of

in uinculis donec probentur quae dixistis utrum falsa an uera Pharaoh surely ye are spies. sint alioquin per salutem Pharaonis exploratores estis 42:17 42:17 And he put them all together into ward three days. tradidit ergo eos custodiae tribus diebus 42:18 die autem 42:18 And Joseph said unto them the third day, This do,

and live; for I fear God:

tertio eductis de carcere ait facite quod dixi et uiuetis Deum enim timeo 42:19 si pacifici estis frater uester unus ligetur 42:19 If ye be true men, let one of your brethren be bound

in the house of your prison: go ye, carry corn for the famine of your houses:

in carcere uos autem abite et ferte frumenta quae emistis in

domos uestras 42:20 et fratrem uestrum minimum ad me 42:20 But bring your youngest brother unto me; so shall your words be verified, and ye shall not die. And they did so.

adducite ut possim uestros probare sermones et non moriamini fecerunt ut dixerat 42:21 et locuti sunt inuicem merito 42:21 And they said one to another, We are verily guilty

concerning our brother, in that we saw the anguish of his soul, when he besought us, and we would not hear; there-

haec patimur quia peccauimus in fratrem nostrum uidentes fore is this distress come upon us. angustiam animae illius cum deprecaretur nos et non audiuimus idcirco uenit super nos ista tribulatio 42:22 e quibus 42:22 And Reuben answered them, saying, Spake I not

unto you, saying, Do not sin against the child; and ye would not hear? therefore, behold, also his blood is re-

unus Ruben ait numquid non dixi uobis nolite peccare in quired. puerum et non audistis me en sanguis eius exquiritur 42:23 42:23 And they knew not that Joseph understood them; for

he spake unto them by an interpreter.

nesciebant autem quod intellegeret Ioseph eo quod per interpretem loquebatur ad eos 42:24 auertitque se parumper et 42:24 And he turned himself about from them, and wept;

and returned to them again, and communed with them, and took from them Simeon, and bound him before their eyes.

fleuit et reuersus locutus est ad eos 42:25 tollens Symeon et

42:25 Then Joseph commanded to fill their sacks with corn, and to restore every man’s money into his sack, and to give them provision for the way: and thus did he unto them.

ligans illis praesentibus iussitque ministris ut implerent saccos eorum tritico et reponerent pecunias singulorum in sacculis suis datis supra cibariis in uia qui fecerunt ita 42:26 at 42:26 And they laded their asses with the corn, and de-

parted thence.

illi portantes frumenta in asinis profecti sunt 42:27 aperto- 42:27 And as one of them opened his sack to give his ass

provender in the inn, he espied his money; for, behold, it was in his sack’s mouth.

que unus sacco ut daret iumento pabulum in diuersorio contemplatus pecuniam in ore sacculi 42:28 dixit fratribus suis 42:28 And he said unto his brethren, My money is re-

stored; and, lo, it is even in my sack: and their heart failed them, and they were afraid, saying one to another, What is

reddita est mihi pecunia en habetur in sacco et obstupefacti this that God hath done unto us? turbatique dixerunt mutuo quidnam est hoc quod fecit no-

Biblia Sacra Vulgata

42:29 And they came unto Jacob their father unto the land bis Deus 42:29 ueneruntque ad Iacob patrem suum in terra of Canaan, and told him all that befell unto them; saying,

Chanaan et narrauerunt ei omnia quae accidissent sibi dicen-

42:30 The man, who is the lord of the land, spake roughly tes 42:30 locutus est nobis dominus terrae dure et putauit

to us, and took us for spies of the country.

42:31 And we said unto him, We are true men; we are no nos exploratores prouinciae 42:31 cui respondimus pacifici

spies: 42:32 We be twelve brethren, sons of our father; one is not, sumus nec ullas molimur insidias 42:32 duodecim fratres and the youngest is this day with our father in the land of Canaan.

uno patre geniti sumus unus non est super minimus cum pa-

42:33 And the man, the lord of the country, said unto us, tre uersatur in terra Chanaan 42:33 qui ait nobis sic probabo

Hereby shall I know that ye are true men; leave one of your brethren here with me, and take food for the famine

of your households, and be gone: quod pacifici sitis fratrem uestrum unum dimittite apud me 42:34 And bring your youngest brother unto me: then et cibaria domibus uestris necessaria sumite et abite 42:34

shall I know that ye are no spies, but that ye are true men: so will I deliver you your brother, and ye shall traffick in

the land. fratremque uestrum minimum adducite ad me ut sciam quod non sitis exploratores et istum qui tenetur in uinculis recipere possitis ac deinceps emendi quae uultis habeatis licentiam

42:35 And it came to pass as they emptied their sacks, that, 42:35 his dictis cum frumenta effunderent singuli reppererunt

behold, every man’s bundle of money was in his sack: and when both they and their father saw the bundles of money,

they were afraid. in ore saccorum ligatas pecunias exterritisque simul omnibus 42:36 And Jacob their father said unto them, Me have ye 42:36 dixit pater Iacob absque liberis me esse fecistis Ioseph

bereaved of my children: Joseph is not, and Simeon is not, and ye will take Benjamin away: all these things are

against me. non est super Symeon tenetur in uinculis Beniamin auferetis 42:37 And Reuben spake unto his father, saying, Slay my in me haec mala omnia reciderunt 42:37 cui respondit Ru-

two sons, if I bring him not to thee: deliver him into my hand, and I will bring him to thee again.

ben duos filios meos interfice si non reduxero illum tibi trade

42:38 And he said, My son shall not go down with you; in manu mea et ego eum restituam 42:38 at ille non descen-

for his brother is dead, and he is left alone: if mischief befall him by the way in the which ye go, then shall ye

bring down my gray hairs with sorrow to the grave. det inquit filius meus uobiscum frater eius mortuus est ipse solus remansit si quid ei aduersi acciderit in terra ad quam pergitis deducetis canos meos cum dolore ad inferos

43:1 And the famine was sore in the land. 43:1 interim fames omnem terram uehementer premebat 43:2 And it came to pass, when they had eaten up the corn 43:2 consumptisque cibis quos ex Aegypto detulerant dixit

which they had brought out of Egypt, their father said unto them, Go again, buy us a little food.

Iacob ad filios suos reuertimini et emite pauxillum escarum

43:3 And Judah spake unto him, saying, The man did 43:3 respondit Iudas denuntiauit nobis uir ille sub testifica-

solemnly protest unto us, saying, Ye shall not see my face, except your brother be with you.

tione iurandi dicens non uidebitis faciem meam nisi fratrem

43:4 If thou wilt send our brother with us, we will go down uestrum minimum adduxeritis uobiscum 43:4 si ergo uis and buy thee food:

mittere eum nobiscum pergemus pariter et ememus tibi ne-

43:5 But if thou wilt not send him, we will not go down: cessaria 43:5 si autem non uis non ibimus uir enim ut saepe

for the man said unto us, Ye shall not see my face, except your brother be with you.

diximus denuntiauit nobis dicens non uidebitis faciem meam

43:6 And Israel said, Wherefore dealt ye so ill with me, as absque fratre uestro minimo 43:6 dixit eis Israhel in meam

to tell the man whether ye had yet a brother?

hoc fecistis miseriam ut indicaretis ei et alium habere uos

Liber Genesis

fratrem 43:7 at illi responderunt interrogauit nos homo per 43:7 And they said, The man asked us straitly of our state,

and of our kindred, saying, Is your father yet alive? have ye another brother? and we told him according to the tenor of these words: could we certainly know that he would say, Bring your brother down?

ordinem nostram progeniem si pater uiueret si haberemus fratrem et nos respondimus ei consequenter iuxta id quod fuerat sciscitatus numquid scire poteramus quod dicturus esset adducite uobiscum fratrem uestrum 43:8 Iudas quoque 43:8 And Judah said unto Israel his father, Send the lad

with me, and we will arise and go; that we may live, and not die, both we, and thou, and also our little ones.

dixit patri suo mitte puerum mecum ut proficiscamur et pos-

simus uiuere ne moriamur nos et paruuli nostri 43:9 ego 43:9 I will be surety for him; of my hand shalt thou require him: if I bring him not unto thee, and set him before thee, then let me bear the blame for ever:

suscipio puerum de manu mea require illum nisi reduxero et tradidero eum tibi ero peccati in te reus omni tempore 43:10 si non intercessisset dilatio iam uice altera uenissemus 43:10 For except we had lingered, surely now we had re-

turned this second time.

43:11 igitur Israhel pater eorum dixit ad eos si sic necesse est 43:11 And their father Israel said unto them, If it must be

so now, do this; take of the best fruits in the land in your vessels, and carry down the man a present, a little balm,

facite quod uultis sumite de optimis terrae fructibus in uasis and a little honey, spices, and myrrh, nuts, and almonds: uestris et deferte uiro munera modicum resinae et mellis et styracis et stactes et terebinthi et amigdalarum 43:12 pecu- 43:12 And take double money in your hand; and the

money that was brought again in the mouth of your sacks, carry it again in your hand; peradventure it was an over-

niamque duplicem ferte uobiscum et illam quam inuenistis sight: in sacculis reportate ne forte errore factum sit 43:13 sed et 43:13 Take also your brother, and arise, go again unto the

man:

fratrem uestrum tollite et ite ad uirum 43:14 Deus autem 43:14 And God Almighty give you mercy before the man,

that he may send away your other brother, and Benjamin. If I be bereaved of my children, I am bereaved.

meus omnipotens faciat uobis eum placabilem et remittat uobiscum fratrem uestrum quem tenet et hunc Beniamin ego autem quasi orbatus absque liberis ero 43:15 tulerunt ergo 43:15 And the men took that present, and they took double

money in their hand and Benjamin; and rose up, and went down to Egypt, and stood before Joseph.

uiri munera et pecuniam duplicem et Beniamin descenderuntque in Aegyptum et steterunt coram Ioseph 43:16 quos 43:16 And when Joseph saw Benjamin with them, he said

to the ruler of his house, Bring these men home, and slay, and make ready; for these men shall dine with me at noon.

cum ille uidisset et Beniamin simul praecepit dispensatori domus suae dicens introduc uiros domum et occide uictimas et instrue conuiuium quoniam mecum sunt comesuri meridie 43:17 fecit ille sicut fuerat imperatum et introduxit ui- 43:17 And the man did as Joseph bade; and the man

brought the men into Joseph’s house.

ros domum 43:18 ibique exterriti dixerunt mutuo propter 43:18 And the men were afraid, because they were brought

into Joseph’s house; and they said, Because of the money that was returned in our sacks at the first time are we brought in; that he may seek occasion against us, and fall upon us, and take us for bondmen, and our asses.

pecuniam quam rettulimus prius in saccis nostris introducti sumus ut deuoluat in nos calumniam et uiolenter subiciat seruituti et nos et asinos nostros 43:19 quam ob rem in ipsis 43:19 And they came near to the steward of Joseph’s

house, and they communed with him at the door of the house,

foribus accedentes ad dispensatorem 43:20 locuti sunt ora-

43:20 And said, O sir, we came indeed down at the first time to buy food:

mus domine ut audias iam ante descendimus ut emeremus escas 43:21 quibus emptis cum uenissemus ad diuersorium 43:21 And it came to pass, when we came to the inn, that

we opened our sacks, and, behold, every man’s money was in the mouth of his sack, our money in full weight: and we have brought it again in our hand.

Biblia Sacra Vulgata

aperuimus sacculos nostros et inuenimus pecuniam in ore

43:22 And other money have we brought down in our saccorum quam nunc eodem pondere reportamus 43:22 sed

hands to buy food: we cannot tell who put our money in our sacks.

et aliud adtulimus argentum ut emamus quae necessaria sunt non est in nostra conscientia quis eam posuerit in marsup-

43:23 And he said, Peace be to you, fear not: your God, piis nostris 43:23 at ille respondit pax uobiscum nolite ti-

and the God of your father, hath given you treasure in your sacks: I had your money. And he brought Simeon out unto

them. mere Deus uester et Deus patris uestri dedit uobis thesauros

in sacculis uestris nam pecuniam quam dedistis mihi pro-

43:24 And the man brought the men into Joseph’s house, batam ego habeo eduxitque ad eos Symeon 43:24 et intro-

and gave them water, and they washed their feet; and he gave their asses provender.

ductis domum adtulit aquam et lauerunt pedes suos deditque

43:25 And they made ready the present against Joseph pabula asinis eorum 43:25 illi uero parabant munera donec

came at noon: for they heard that they should eat bread there.

ingrederetur Ioseph meridie audierant enim quod ibi come-

43:26 And when Joseph came home, they brought him the suri essent panem 43:26 igitur ingressus est Ioseph domum

present which was in their hand into the house, and bowed themselves to him to the earth.

suam obtuleruntque ei munera tenentes in manibus et ad-

43:27 And he asked them of their welfare, and said, Is orauerunt proni in terram 43:27 at ille clementer resalutatis

your father well, the old man of whom ye spake? Is he yet alive?

eis interrogauit dicens saluusne est pater uester senex de quo

43:28 And they answered, Thy servant our father is in dixeratis mihi adhuc uiuit 43:28 qui responderunt sospes est

good health, he is yet alive. And they bowed down their heads, and made obeisance.

seruus tuus pater noster adhuc uiuit et incuruati adorauerunt

43:29 And he lifted up his eyes, and saw his brother Ben- eum 43:29 adtollens autem oculos Ioseph uidit Beniamin

jamin, his mother’s son, and said, Is this your younger brother, of whom ye spake unto me? And he said, God be

gracious unto thee, my son. fratrem suum uterinum et ait iste est frater uester paruulus de

quo dixeratis mihi et rursum Deus inquit misereatur tui fili

43:30 And Joseph made haste; for his bowels did yearn mi 43:30 festinauitque quia commota fuerant uiscera eius

upon his brother: and he sought where to weep; and he entered into his chamber, and wept there.

super fratre suo et erumpebant lacrimae et introiens cubi-

43:31 And he washed his face, and went out, and refrained culum fleuit 43:31 rursusque lota facie egressus continuit

himself, and said, Set on bread.

43:32 And they set on for him by himself, and for them se et ait ponite panes 43:32 quibus adpositis seorsum Io-

by themselves, and for the Egyptians, which did eat with him, by themselves: because the Egyptians might not eat bread with the Hebrews; for that is an abomination unto the Egyptians.

seph et seorsum fratribus Aegyptiis quoque qui uescebantur simul seorsum inlicitum est enim Aegyptiis comedere cum Hebraeis et profanum putant huiuscemodi conuiuium

43:33 And they sat before him, the firstborn according to 43:33 sederunt coram eo primogenitus iuxta primogenita sua

his birthright, and the youngest according to his youth: and the men marvelled one at another.

et minimus iuxta aetatem suam et mirabantur nimis 43:34

43:34 And he took and sent messes unto them from before him: but Benjamin’s mess was five times so much as any of their’s. And they drank, and were merry with him.

sumptis partibus quas ab eo acceperant maiorque pars uenit Beniamin ita ut quinque partibus excederet biberuntque et inebriati sunt cum eo

44:1 And he commanded the steward of his house, saying, 44:1 praecepit autem Ioseph dispensatori domus suae di-

Fill the men’s sacks with food, as much as they can carry, and put every man’s money in his sack’s mouth.

Liber Genesis

cens imple saccos eorum frumento quantum possunt capere et pone pecuniam singulorum in summitate sacci 44:2 scy- 44:2 And put my cup, the silver cup, in the sack’s mouth of

the youngest, and his corn money. And he did according to the word that Joseph had spoken.

phum autem meum argenteum et pretium quod dedit tritici pone in ore sacci iunioris factumque est ita 44:3 et orto 44:3 As soon as the morning was light, the men were sent

away, they and their asses.

mane dimissi sunt cum asinis suis 44:4 iamque urbem exie- 44:4 And when they were gone out of the city, and not yet

far off, Joseph said unto his steward, Up, follow after the men; and when thou dost overtake them, say unto them,

rant et processerant paululum tum Ioseph arcessito dispensa- Wherefore have ye rewarded evil for good? tore domus surge inquit persequere uiros et adprehensis dicito quare reddidistis malum pro bono 44:5 scyphum quem 44:5 Is not this it in which my lord drinketh, and whereby

indeed he divineth? ye have done evil in so doing.

furati estis ipse est in quo bibit dominus meus et in quo augurari solet pessimam rem fecistis 44:6 fecit ille ut iusserat et 44:6 And he overtook them, and he spake unto them these

same words.

adprehensis per ordinem locutus est 44:7 qui responderunt 44:7 And they said unto him, Wherefore saith my lord

these words? God forbid that thy servants should do according to this thing:

quare sic loquitur dominus noster ut serui tui tantum flagitii commiserint 44:8 pecuniam quam inuenimus in summi- 44:8 Behold, the money, which we found in our sacks’

mouths, we brought again unto thee out of the land of Canaan: how then should we steal out of thy lord’s house

tate saccorum reportauimus ad te de terra Chanaan et quo- silver or gold? modo consequens est ut furati simus de domo domini tui aurum uel argentum 44:9 apud quemcumque fuerit inuentum 44:9 With whomsoever of thy servants it be found, both

let him die, and we also will be my lord’s bondmen.

seruorum tuorum quod quaeris moriatur et nos serui erimus domini nostri 44:10 qui dixit fiat iuxta uestram sententiam 44:10 And he said, Now also let it be according unto your

words: he with whom it is found shall be my servant; and ye shall be blameless.

apud quem fuerit inuentum ipse sit seruus meus uos autem eritis innoxii 44:11 itaque festinato deponentes in terram 44:11 Then they speedily took down every man his sack

to the ground, and opened every man his sack.

saccos aperuerunt singuli 44:12 quos scrutatus incipiens a 44:12 And he searched, and began at the eldest, and left at

the youngest: and the cup was found in Benjamin’s sack.

maiore usque ad minimum inuenit scyphum in sacco Beniamin 44:13 at illi scissis uestibus oneratisque rursum asinis 44:13 Then they rent their clothes, and laded every man

his ass, and returned to the city.

reuersi sunt in oppidum 44:14 primusque Iudas cum frat- 44:14 And Judah and his brethren came to Joseph’s house;

for he was yet there: and they fell before him on the ground.

ribus ingressus est ad Ioseph necdum enim de loco abierat omnesque ante eum in terra pariter corruerunt 44:15 qui- 44:15 And Joseph said unto them, What deed is this that

ye have done? wot ye not that such a man as I can certainly divine?

bus ille ait cur sic agere uoluistis an ignoratis quod non sit similis mei in augurandi scientia 44:16 cui Iudas quid res- 44:16 And Judah said, What shall we say unto my lord?

what shall we speak? or how shall we clear ourselves? God hath found out the iniquity of thy servants: behold, we are my lord’s servants, both we, and he also with whom the cup is found.

pondebimus inquit domino meo uel quid loquemur aut iusti poterimus obtendere Deus inuenit iniquitatem seruorum tuorum en omnes serui sumus domini mei et nos et apud quem inuentus est scyphus 44:17 respondit Ioseph absit a me ut 44:17 And he said, God forbid that I should do so: but

the man in whose hand the cup is found, he shall be my servant; and as for you, get you up in peace unto your

sic agam qui furatus est scyphum ipse sit seruus meus uos father.

Biblia Sacra Vulgata

44:18 Then Judah came near unto him, and said, Oh my autem abite liberi ad patrem uestrum 44:18 accedens pro-

lord, let thy servant, I pray thee, speak a word in my lord’s ears, and let not thine anger burn against thy servant: for

thou art even as Pharaoh. pius Iudas confidenter ait oro domine mi loquatur seruus

tuus uerbum in auribus tuis et ne irascaris famulo tuo tu

44:19 My lord asked his servants, saying, Have ye a father, es enim post Pharaonem 44:19 dominus meus interrogasti or a brother?

44:20 And we said unto my lord, We have a father, an old prius seruos tuos habetis patrem aut fratrem 44:20 et nos

man, and a child of his old age, a little one; and his brother is dead, and he alone is left of his mother, and his father

loveth him. respondimus tibi domino meo est nobis pater senex et puer paruulus qui in senecta illius natus est cuius uterinus frater est mortuus et ipsum solum habet mater sua pater uero te-

44:21 And thou saidst unto thy servants, Bring him down nere diligit eum 44:21 dixistique seruis tuis adducite eum

unto me, that I may set mine eyes upon him. 44:22 And we said unto my lord, The lad cannot leave his ad me et ponam oculos meos super illum 44:22 suggessifather: for if he should leave his father, his father would die.

mus domino meo non potest puer relinquere patrem suum

44:23 And thou saidst unto thy servants, Except your si enim illum dimiserit morietur 44:23 et dixisti seruis tuis

youngest brother come down with you, ye shall see my face no more.

nisi uenerit frater uester minimus uobiscum non uidebitis

44:24 And it came to pass when we came up unto thy ser- amplius faciem meam 44:24 cum ergo ascendissemus ad

vant my father, we told him the words of my lord.

famulum tuum patrem nostrum narrauimus ei omnia quae

44:25 And our father said, Go again, and buy us a little locutus est dominus meus 44:25 et dixit pater noster reuer-

food.

44:26 And we said, We cannot go down: if our youngest timini et emite nobis parum tritici 44:26 cui diximus ire non

brother be with us, then will we go down: for we may not see the man’s face, except our youngest brother be with

us. possumus si frater noster minimus descendet nobiscum pro-

ficiscemur simul alioquin illo absente non audemus uidere

44:27 And thy servant my father said unto us, Ye know faciem uiri 44:27 atque ille respondit uos scitis quod duos

that my wife bare me two sons:

44:28 And the one went out from me, and I said, Surely genuerit mihi uxor mea 44:28 egressus est unus et dixistis

he is torn in pieces; and I saw him not since:

44:29 And if ye take this also from me, and mischief befall bestia deuorauit eum et hucusque non conparet 44:29 si tu-

him, ye shall bring down my gray hairs with sorrow to the grave.

leritis et istum et aliquid ei in uia contigerit deducetis canos

44:30 Now therefore when I come to thy servant my father, meos cum maerore ad inferos 44:30 igitur si intrauero ad and the lad be not with us; seeing that his life is bound up in the lad’s life;

seruum tuum patrem nostrum et puer defuerit cum anima il-

44:31 It shall come to pass, when he seeth that the lad is lius ex huius anima pendeat 44:31 uideritque eum non esse

not with us, that he will die: and thy servants shall bring down the gray hairs of thy servant our father with sorrow

to the grave. nobiscum morietur et deducent famuli tui canos eius cum 44:32 For thy servant became surety for the lad unto my dolore ad inferos 44:32 ego proprie seruus tuus qui in meam

father, saying, If I bring him not unto thee, then I shall bear the blame to my father for ever.

hunc recepi fidem et spopondi dicens nisi reduxero eum pec-

44:33 Now therefore, I pray thee, let thy servant abide in- cati reus ero in patrem meum omni tempore 44:33 manebo stead of the lad a bondman to my lord; and let the lad go up with his brethren.

itaque seruus tuus pro puero in ministerium domini mei et

44:34 For how shall I go up to my father, and the lad be not puer ascendat cum fratribus suis 44:34 non enim possum with me? lest peradventure I see the evil that shall come on my father.

rediie ad patrem absente puero ne calamitatis quae oppres-

Liber Genesis

sura est patrem meum testis adsistam

45:1 non se poterat ultra cohibere Ioseph multis coram ad- 45:1 Then Joseph could not refrain himself before all them

that stood by him; and he cried, Cause every man to go out from me. And there stood no man with him, while Joseph

stantibus unde praecepit ut egrederentur cuncti foras et nul- made himself known unto his brethren. lus interesset alienus agnitioni mutuae 45:2 eleuauitque uo- 45:2 And he wept aloud: and the Egyptians and the house

of Pharaoh heard.

cem cum fletu quam audierunt Aegyptii omnisque domus Pharaonis 45:3 et dixit fratribus suis ego sum Ioseph ad- 45:3 And Joseph said unto his brethren, I am Joseph; doth

my father yet live? And his brethren could not answer him; for they were troubled at his presence.

huc pater meus uiuit nec poterant respondere fratres nimio timore perterriti 45:4 ad quos ille clementer accedite inquit 45:4 And Joseph said unto his brethren, Come near to me,

I pray you. And they came near. And he said, I am Joseph your brother, whom ye sold into Egypt.

ad me et cum accessissent prope ego sum ait Ioseph frater uester quem uendidistis in Aegypto 45:5 nolite pauere nec 45:5 Now therefore be not grieved, nor angry with your-

selves, that ye sold me hither: for God did send me before you to preserve life.

uobis durum esse uideatur quod uendidistis me in his regionibus pro salute enim uestra misit me Deus ante uos in Aegyptum 45:6 biennium est quod fames esse coepit in terra 45:6 For these two years hath the famine been in the land:

and yet there are five years, in the which there shall neither be earing nor harvest.

et adhuc quinque anni restant quibus nec arari poterit nec meti 45:7 praemisitque me Deus ut reseruemini super ter- 45:7 And God sent me before you to preserve you a pos-

terity in the earth, and to save your lives by a great deliverance.

ram et escas ad uiuendum habere possitis 45:8 non uestro

45:8 So now it was not you that sent me hither, but God: and he hath made me a father to Pharaoh, and lord of all his house, and a ruler throughout all the land of Egypt.

consilio sed Dei huc uoluntate missus sum qui fecit me quasi patrem Pharaonis et dominum uniuersae domus eius ac principem in omni terra Aegypti 45:9 festinate et ascendite ad 45:9 Haste ye, and go up to my father, and say unto him,

Thus saith thy son Joseph, God hath made me lord of all Egypt: come down unto me, tarry not:

patrem meum et dicetis ei haec mandat filius tuus Ioseph Deus me fecit dominum uniuersae terrae Aegypti descende ad me ne moreris 45:10 et habita in terra Gessen erisque 45:10 And thou shalt dwell in the land of Goshen, and

thou shalt be near unto me, thou, and thy children, and thy children’s children, and thy flocks, and thy herds, and all

iuxta me tu et filii tui et filii filiorum tuorum oues tuae et ar- that thou hast: menta tua et uniuersa quae possides 45:11 ibique te pascam 45:11 And there will I nourish thee; for yet there are five

years of famine; lest thou, and thy household, and all that thou hast, come to poverty.

adhuc enim quinque anni residui sunt famis ne et tu pereas et domus tua et omnia quae possides 45:12 en oculi uestri 45:12 And, behold, your eyes see, and the eyes of my

brother Benjamin, that it is my mouth that speaketh unto you.

et oculi fratris mei Beniamin uident quod os meum loquatur

ad uos 45:13 nuntiate patri meo uniuersam gloriam meam 45:13 And ye shall tell my father of all my glory in Egypt, and of all that ye have seen; and ye shall haste and bring down my father hither.

et cuncta quae uidistis in Aegypto festinate et adducite eum ad me 45:14 cumque amplexatus recidisset in collum Be- 45:14 And he fell upon his brother Benjamin’s neck, and

wept; and Benjamin wept upon his neck.

niamin fratris sui fleuit illo quoque flente similiter super collum eius 45:15 osculatusque est Ioseph omnes fratres suos 45:15 Moreover he kissed all his brethren, and wept upon

them: and after that his brethren talked with him.

et plorauit super singulos post quae ausi sunt loqui ad eum

Biblia Sacra Vulgata

45:16 And the fame thereof was heard in Pharaoh’s 45:16 auditumque est et celebri sermone uulgatum in aula re-

house, saying, Joseph’s brethren are come: and it pleased Pharaoh well, and his servants.

gis uenerunt fratres Ioseph et gauisus est Pharao atque om-

45:17 And Pharaoh said unto Joseph, Say unto thy nis familia eius 45:17 dixitque ad Ioseph ut imperaret fra-

brethren, This do ye; lade your beasts, and go, get you unto the land of Canaan;

tribus suis dicens onerantes iumenta ite in terram Chanaan

45:18 And take your father and your households, and 45:18 et tollite inde patrem uestrum et cognationem et uenite

come unto me: and I will give you the good of the land of Egypt, and ye shall eat the fat of the land.

ad me et ego dabo uobis omnia bona Aegypti ut comedatis

45:19 Now thou art commanded, this do ye; take you wag- medullam terrae 45:19 praecipe etiam ut tollant plaustra de

ons out of the land of Egypt for your little ones, and for your wives, and bring your father, and come.

terra Aegypti ad subuectionem paruulorum suorum et coniugum ac dicito tollite patrem uestrum et properate quanto-

45:20 Also regard not your stuff; for the good of all the cius uenientes 45:20 ne dimittatis quicquam de supellectili

land of Egypt is your’s.

45:21 And the children of Israel did so: and Joseph gave uestra quia omnes opes Aegypti uestrae erunt 45:21 fece-

them wagons, according to the commandment of Pharaoh, and gave them provision for the way.

runt filii Israhel ut eis mandatum fuerat quibus dedit Ioseph plaustra secundum Pharaonis imperium et cibaria in itinere

45:22 To all of them he gave each man changes of raiment; 45:22 singulisque proferri iussit binas stolas Beniamin uero

but to Benjamin he gave three hundred pieces of silver, and five changes of raiment.

dedit trecentos argenteos cum quinque stolis optimis 45:23

45:23 And to his father he sent after this manner; ten asses laden with the good things of Egypt, and ten she asses laden with corn and bread and meat for his father by the way.

tantundem pecuniae et uestium mittens patri suo addens eis asinos decem qui subueherent ex omnibus diuitiis Aegypti et

45:24 So he sent his brethren away, and they departed: and totidem asinas triticum in itinere panesque portantes 45:24

he said unto them, See that ye fall not out by the way.

dimisit ergo fratres suos et proficiscentibus ait ne irascamini

45:25 And they went up out of Egypt, and came into the in uia 45:25 qui ascendentes ex Aegypto uenerunt in terram

land of Canaan unto Jacob their father,

45:26 And told him, saying, Joseph is yet alive, and he Chanaan ad patrem suum Iacob 45:26 et nuntiauerunt ei di-

is governor over all the land of Egypt. And Jacob’s heart fainted, for he believed them not.

centes Ioseph uiuit et ipse dominatur in omni terra Aegypti quo audito quasi de graui somno euigilans tamen non crede-

45:27 And they told him all the words of Joseph, which he bat eis 45:27 illi contra referebant omnem ordinem rei cum-

had said unto them: and when he saw the wagons which Joseph had sent to carry him, the spirit of Jacob their father

revived: que uidisset plaustra et uniuersa quae miserat reuixit spiritus 45:28 And Israel said, It is enough; Joseph my son is yet eius 45:28 et ait sufficit mihi si adhuc Ioseph filius meus

alive: I will go and see him before I die.

uiuit uadam et uidebo illum antequam moriar

46:1 And Israel took his journey with all that he had, and 46:1 profectusque Israhel cum omnibus quae habebat uecame to Beersheba, and offered sacrifices unto the God of his father Isaac.

nit ad puteum Iuramenti et mactatis ibi uictimis Deo patris

46:2 And God spake unto Israel in the visions of the night, sui Isaac 46:2 audiuit eum per uisionem nocte uocantem se and said, Jacob, Jacob. And he said, Here am I.

46:3 And he said, I am God, the God of thy father: fear et dicentem sibi Iacob Iacob cui respondit ecce adsum 46:3

not to go down into Egypt; for I will there make of thee a great nation:

ait illi Deus ego sum Fortissimus Deus patris tui noli timere et descende in Aegyptum quia in gentem magnam faciam te

Liber Genesis

ibi 46:4 ego descendam tecum illuc et ego inde adducam te 46:4 I will go down with thee into Egypt; and I will also

surely bring thee up again: and Joseph shall put his hand upon thine eyes.

reuertentem Ioseph quoque ponet manum suam super oculos tuos 46:5 surrexit Iacob a puteo Iuramenti tuleruntque eum 46:5 And Jacob rose up from Beersheba: and the sons of

Israel carried Jacob their father, and their little ones, and their wives, in the wagons which Pharaoh had sent to carry

filii cum paruulis et uxoribus suis in plaustris quae miserat him. Pharao ad portandum senem 46:6 et omnia quae possederat 46:6 And they took their cattle, and their goods, which

they had gotten in the land of Canaan, and came into Egypt, Jacob, and all his seed with him:

in terra Chanaan uenitque in Aegyptum cum omni semine suo 46:7 filii eius et nepotes filiae et cuncta simul proge- 46:7 His sons, and his sons’ sons with him, his daughters,

and his sons’ daughters, and all his seed brought he with him into Egypt.

nies 46:8 haec sunt autem nomina filiorum Israhel qui in-

46:8 And these are the names of the children of Israel, which came into Egypt, Jacob and his sons: Reuben, Jacob’s firstborn.

gressi sunt in Aegyptum ipse cum liberis suis primogenitus Ruben 46:9 filii Ruben Enoch et Phallu et Esrom et Charmi 46:9 And the sons of Reuben; Hanoch, and Phallu, and

Hezron, and Carmi.

46:10 filii Symeon Iemuhel et Iamin et Ahod et Iachin et Sa- 46:10 And the sons of Simeon; Jemuel, and Jamin, and

Ohad, and Jachin, and Zohar, and Shaul the son of a Canaanitish woman.

her et Saul filius Chananitidis 46:11 filii Leui Gerson Caath

46:11 And the sons of Levi; Gershon, Kohath, and Merari.

et Merari 46:12 filii Iuda Her et Onan et Sela et Phares et 46:12 And the sons of Judah; Er, and Onan, and Shelah,

and Pharez, and Zarah: but Er and Onan died in the land of Canaan. And the sons of Pharez were Hezron and Hamul.

Zara mortui sunt autem Her et Onan in terra Chanaan natique sunt filii Phares Esrom et Amul 46:13 filii Isachar Thola et 46:13 And the sons of Issachar; Tola, and Phuvah, and

Job, and Shimron.

Phua et Iob et Semron 46:14 filii Zabulon Sared et Helon 46:14 And the sons of Zebulun; Sered, and Elon, and

Jahleel.

et Iahelel 46:15 hii filii Liae quos genuit in Mesopotamiam 46:15 These be the sons of Leah, which she bare unto Ja-

cob in Padanaram, with his daughter Dinah: all the souls of his sons and his daughters were thirty and three.

Syriae cum Dina filia sua omnes animae filiorum eius et filiarum triginta tres 46:16 filii Gad Sephion et Haggi Suni et 46:16 And the sons of Gad; Ziphion, and Haggi, Shuni,

and Ezbon, Eri, and Arodi, and Areli.

Esebon Heri et Arodi et Areli 46:17 filii Aser Iamne et Iesua 46:17 And the sons of Asher; Jimnah, and Ishuah, and

Isui, and Beriah, and Serah their sister: and the sons of Beriah; Heber, and Malchiel.

et Iesui et Beria Sara quoque soror eorum filii Beria Heber et Melchihel 46:18 hii filii Zelphae quam dedit Laban Liae 46:18 These are the sons of Zilpah, whom Laban gave to

Leah his daughter, and these she bare unto Jacob, even sixteen souls.

filiae suae et hos genuit Iacob sedecim animas 46:19 filii

46:19 The sons of Rachel Jacob’s wife; Joseph, and Benjamin.

Rahel uxoris Iacob Ioseph et Beniamin 46:20 natique sunt 46:20 And unto Joseph in the land of Egypt were born

Manasseh and Ephraim, which Asenath the daughter of Potipherah priest of On bare unto him.

Ioseph filii in terra Aegypti quos genuit ei Aseneth filia Putiphare sacerdotis Heliopoleos Manasses et Ephraim 46:21 46:21 And the sons of Benjamin were Belah, and Becher,

and Ashbel, Gera, and Naaman, Ehi, and Rosh, Muppim, and Huppim, and Ard.

filii Beniamin Bela et Bechor et Asbel Gera et Naaman et Ehi et Ros Mophim et Opphim et Ared 46:22 hii filii Rahel 46:22 These are the sons of Rachel, which were born to

Jacob: all the souls were fourteen.

quos genuit Iacob omnes animae quattuordecim 46:23 filii 46:23 And the sons of Dan; Hushim. Dan Vsim 46:24 filii Nepthalim Iasihel et Guni et Hieser 46:24 And the sons of Naphtali; Jahzeel, and Guni, and

Jezer, and Shillem.

et Sallem 46:25 hii filii Balae quam dedit Laban Raheli fi- 46:25 These are the sons of Bilhah, which Laban gave unto

Rachel his daughter, and she bare these unto Jacob: all the souls were seven.

liae suae et hos genuit Iacob omnes animae septem 46:26

46:26 All the souls that came with Jacob into Egypt, which came out of his loins, besides Jacob’s sons’ wives, all the souls were threescore and six;

cunctae animae quae ingressae sunt cum Iacob in Aegyptum

Biblia Sacra Vulgata

et egressae de femore illius absque uxoribus filiorum sexa-

46:27 And the sons of Joseph, which were born him in ginta sex 46:27 filii autem Ioseph qui nati sunt ei in terra

Egypt, were two souls: all the souls of the house of Jacob, which came into Egypt, were threescore and ten.

Aegypti animae duae omnis anima domus Iacob quae in-

46:28 And he sent Judah before him unto Joseph, to di- gressa est Aegyptum fuere septuaginta 46:28 misit autem

rect his face unto Goshen; and they came into the land of Goshen.

Iudam ante se ad Ioseph ut nuntiaret ei et ille occurreret in

46:29 And Joseph made ready his chariot, and went up to Gessen 46:29 quo cum peruenisset iuncto Ioseph curru suo

meet Israel his father, to Goshen, and presented himself unto him; and he fell on his neck, and wept on his neck a

good while. ascendit obuiam patri ad eundem locum uidensque eum in-46:30 And Israel said unto Joseph, Now let me die, since ruit super collum eius et inter amplexus fleuit 46:30 dixit-

I have seen thy face, because thou art yet alive.

que pater ad Ioseph iam laetus moriar quia uidi faciem tuam

46:31 And Joseph said unto his brethren, and unto his fa- et superstitem te relinquo 46:31 et ille locutus est ad fratres

ther’s house, I will go up, and shew Pharaoh, and say unto him, My brethren, and my father’s house, which were in

the land of Canaan, are come unto me; et ad omnem domum patris sui ascendam et nuntiabo Pha-

raoni dicamque ei fratres mei et domus patris mei qui erant

46:32 And the men are shepherds, for their trade hath been in terra Chanaan uenerunt ad me 46:32 et sunt uiri pastores

to feed cattle; and they have brought their flocks, and their herds, and all that they have.

ouium curamque habent alendorum gregum pecora sua et armenta et omnia quae habere potuerunt adduxerunt secum

46:33 And it shall come to pass, when Pharaoh shall call 46:33 cumque uocauerit uos et dixerit quod est opus uestrum

you, and shall say, What is your occupation?

46:34 That ye shall say, Thy servants’ trade hath been 46:34 respondebitis uiri pastores sumus serui tui ab infantia

about cattle from our youth even until now, both we, and also our fathers: that ye may dwell in the land of Goshen;

for every shepherd is an abomination unto the Egyptians. nostra usque in praesens et nos et patres nostri haec autem dicetis ut habitare possitis in terra Gessen quia detestantur Aegyptii omnes pastores ouium

47:1 Then Joseph came and told Pharaoh, and said, My 47:1 ingressus ergo Ioseph nuntiauit Pharaoni dicens pa-

father and my brethren, and their flocks, and their herds, and all that they have, are come out of the land of Canaan;

and, behold, they are in the land of Goshen. ter meus et fratres oues eorum et armenta et cuncta quae

possident uenerunt de terra Chanaan et ecce consistunt in

47:2 And he took some of his brethren, even five men, and terra Gessen 47:2 extremos quoque fratrum suorum quin-

presented them unto Pharaoh.

47:3 And Pharaoh said unto his brethren, What is your que uiros statuit coram rege 47:3 quos ille interrogauit quid

occupation? And they said unto Pharaoh, Thy servants are shepherds, both we, and also our fathers.

habetis operis responderunt pastores ouium sumus serui tui

47:4 They said morever unto Pharaoh, For to sojourn in et nos et patres nostri 47:4 ad peregrinandum in terra tua

the land are we come; for thy servants have no pasture for their flocks; for the famine is sore in the land of Canaan: now therefore, we pray thee, let thy servants dwell in the land of Goshen.

uenimus quoniam non est herba gregibus seruorum tuorum ingrauescente fame in regione Chanaan petimusque ut esse

47:5 And Pharaoh spake unto Joseph, saying, Thy father nos iubeas seruos tuos in terra Gessen 47:5 dixit itaque rex

and thy brethren are come unto thee:

47:6 The land of Egypt is before thee; in the best of the ad Ioseph pater tuus et fratres tui uenerunt ad te 47:6 terra

land make thy father and brethren to dwell; in the land of Goshen let them dwell: and if thou knowest any men of activity among them, then make them rulers over my cattle.

Aegypti in conspectu tuo est in optimo loco fac habitare eos et trade eis terram Gessen quod si nosti esse in eis uiros indu-

Liber Genesis

strios constitue illos magistros pecorum meorum 47:7 post 47:7 And Joseph brought in Jacob his father, and set him

before Pharaoh: and Jacob blessed Pharaoh.

haec introduxit Ioseph patrem suum ad regem et statuit eum coram eo qui benedicens illi 47:8 et interrogatus ab eo quot 47:8 And Pharaoh said unto Jacob, How old art thou? sunt dies annorum uitae tuae 47:9 respondit dies peregri- 47:9 And Jacob said unto Pharaoh, The days of the years

of my pilgrimage are an hundred and thirty years: few and evil have the days of the years of my life been, and have not attained unto the days of the years of the life of my fathers in the days of their pilgrimage.

nationis uitae meae centum triginta annorum sunt parui et mali et non peruenerunt usque ad dies patrum meorum quibus peregrinati sunt 47:10 et benedicto rege egressus est fo- 47:10 And Jacob blessed Pharaoh, and went out from be-

fore Pharaoh.

ras 47:11 Ioseph uero patri et fratribus suis dedit posses- 47:11 And Joseph placed his father and his brethren, and

gave them a possession in the land of Egypt, in the best of the land, in the land of Rameses, as Pharaoh had com-

sionem in Aegypto in optimo loco terrae solo Ramesses ut manded.

praeceperat Pharao 47:12 et alebat eos omnemque domum 47:12 And Joseph nourished his father, and his brethren, and all his father’s household, with bread, according to their families.

patris sui praebens cibaria singulis 47:13 in toto enim orbe

47:13 And there was no bread in all the land; for the famine was very sore, so that the land of Egypt and all the land of Canaan fainted by reason of the famine.

panis deerat et oppresserat fames terram maxime Aegypti et Chanaan 47:14 e quibus omnem pecuniam congregauit pro 47:14 And Joseph gathered up all the money that was

found in the land of Egypt, and in the land of Canaan, for the corn which they bought: and Joseph brought the

uenditione frumenti et intulit eam in aerarium regis 47:15 money into Pharaoh’s house.

47:15 And when money failed in the land of Egypt, and in the land of Canaan, all the Egyptians came unto Joseph, and said, Give us bread: for why should we die in thy presence? for the money faileth.

cumque defecisset emptoris pretium uenit cuncta Aegyptus ad Ioseph dicens da nobis panes quare morimur coram te deficiente pecunia 47:16 quibus ille respondit adducite pe- 47:16 And Joseph said, Give your cattle; and I will give

you for your cattle, if money fail.

cora uestra et dabo uobis pro eis cibos si pretium non habetis 47:17 quae cum adduxissent dedit eis alimenta pro equis 47:17 And they brought their cattle unto Joseph: and

Joseph gave them bread in exchange for horses, and for the flocks, and for the cattle of the herds, and for the asses:

et ouibus et bubus et asinis sustentauitque eos illo anno pro and he fed them with bread for all their cattle for that year. commutatione pecorum 47:18 ueneruntque anno secundo et 47:18 When that year was ended, they came unto him the

second year, and said unto him, We will not hide it from my lord, how that our money is spent; my lord also hath our herds of cattle; there is not ought left in the sight of my lord, but our bodies, and our lands:

dixerunt ei non celamus dominum nostrum quod deficiente pecunia pecora simul defecerint nec clam te est quod absque corporibus et terra nihil habeamus 47:19 cur ergo mo- 47:19 Wherefore shall we die before thine eyes, both we

and our land? buy us and our land for bread, and we and our land will be servants unto Pharaoh: and give us seed,

rimur te uidente et nos et terra nostra tui erimus eme nos that we may live, and not die, that the land be not desolate.

in seruitutem regiam et praebe semina ne pereunte cultore rediiatur terra in solitudinem 47:20 emit igitur Ioseph om- 47:20 And Joseph bought all the land of Egypt for

Pharaoh; for the Egyptians sold every man his field, because the famine prevailed over them: so the land became

nem terram Aegypti uendentibus singulis possessiones suas Pharaoh’s. prae magnitudine famis subiecitque eam Pharaoni 47:21 et 47:21 And as for the people, he removed them to cities

from one end of the borders of Egypt even to the other end thereof.

cunctos populos eius a nouissimis terminis Aegypti usque ad extremos fines eius 47:22 praeter terram sacerdotum quae 47:22 Only the land of the priests bought he not; for the

priests had a portion assigned them of Pharaoh, and did eat their portion which Pharaoh gave them: wherefore they

a rege tradita fuerat eis quibus et statuta cibaria ex horreis sold not their lands. publicis praebebantur et idcirco non sunt conpulsi uendere

Biblia Sacra Vulgata

47:23 Then Joseph said unto the people, Behold, I have possessiones suas 47:23 dixit ergo Ioseph ad populos en ut

bought you this day and your land for Pharaoh: lo, here is seed for you, and ye shall sow the land.

cernitis et uos et terram uestram Pharao possidet accipite se-

47:24 And it shall come to pass in the increase, that ye mina et serite agros 47:24 ut fruges habere possitis quintam

shall give the fifth part unto Pharaoh, and four parts shall be your own, for seed of the field, and for your food, and for them of your households, and for food for your little ones.

partem regi dabitis quattuor reliquas permitto uobis in se-

47:25 And they said, Thou hast saved our lives: let us find mentem et in cibos famulis et liberis uestris 47:25 qui res-

grace in the sight of my lord, and we will be Pharaoh’s servants.

ponderunt salus nostra in manu tua est respiciat nos tantum

47:26 And Joseph made it a law over the land of Egypt dominus noster et laeti seruiemus regi 47:26 ex eo tempore

unto this day, that Pharaoh should have the fifth part, except the land of the priests only, which became not

Pharaoh’s. usque in praesentem diem in uniuersa terra Aegypti regibus

quinta pars soluitur et factum est quasi in legem absque terra

47:27 And Israel dwelt in the land of Egypt, in the country sacerdotali quae libera ab hac condicione fuit 47:27 habita-

of Goshen; and they had possessions therein, and grew, and multiplied exceedingly.

uit ergo Israhel in Aegypto id est in terra Gessen et posse-

47:28 And Jacob lived in the land of Egypt seventeen dit eam auctusque est et multiplicatus nimis 47:28 et uixit

years: so the whole age of Jacob was an hundred forty and seven years.

in ea decem et septem annis factique sunt omnes dies uitae

47:29 And the time drew nigh that Israel must die: and illius centum quadraginta septem annorum 47:29 cumque

he called his son Joseph, and said unto him, If now I have found grace in thy sight, put, I pray thee, thy hand under my thigh, and deal kindly and truly with me; bury me not, I pray thee, in Egypt:

adpropinquare cerneret mortis diem uocauit filium suum Ioseph et dixit ad eum si inueni gratiam in conspectu tuo pone manum sub femore meo et facies mihi misericordiam et ue-

47:30 But I will lie with my fathers, and thou shalt carry ritatem ut non sepelias me in Aegypto 47:30 sed dormiam me out of Egypt, and bury me in their buryingplace. And he said, I will do as thou hast said.

cum patribus meis et auferas me de hac terra condasque in sepulchro maiorum cui respondit Ioseph ego faciam quod

47:31 And he said, Swear unto me. And he sware unto iussisti 47:31 et ille iura ergo inquit mihi quo iurante adora-

him. And Israel bowed himself upon the bed’s head.

uit Israhel Deum conuersus ad lectuli caput

48:1 And it came to pass after these things, that one told 48:1 his ita transactis nuntiatum est Ioseph quod aegrotaret

Joseph, Behold, thy father is sick: and he took with him his two sons, Manasseh and Ephraim.

pater eius qui adsumptis duobus filiis Manasse et Ephraim

48:2 And one told Jacob, and said, Behold, thy son Joseph ire perrexit 48:2 dictumque est seni ecce filius tuus Ioseph

cometh unto thee: and Israel strengthened himself, and sat upon the bed.

uenit ad te qui confortatus sedit in lectulo 48:3 et ingresso

48:3 And Jacob said unto Joseph, God Almighty appeared unto me at Luz in the land of Canaan, and blessed me,

ad se ait Deus omnipotens apparuit mihi in Luza quae est in

48:4 And said unto me, Behold, I will make thee fruitful, terra Chanaan benedixitque mihi 48:4 et ait ego te augebo

and multiply thee, and I will make of thee a multitude of people; and will give this land to thy seed after thee for an

everlasting possession. et multiplicabo et faciam in turbas populorum daboque tibi

terram hanc et semini tuo post te in possessionem sempiter-

48:5 And now thy two sons, Ephraim and Manasseh, nam 48:5 duo igitur filii tui qui nati sunt tibi in terra Ae-

which were born unto thee in the land of Egypt before I came unto thee into Egypt, are mine; as Reuben and

Simeon, they shall be mine. gypti antequam huc uenirem ad te mei erunt Ephraim et Ma-48:6 And thy issue, which thou begettest after them, nasses sicut Ruben et Symeon reputabuntur mihi 48:6 reli-

shall be thine, and shall be called after the name of their brethren in their inheritance.

Liber Genesis

quos autem quos genueris post eos tui erunt et nomine fratrum suorum uocabuntur in possessionibus suis 48:7 mihi 48:7 And as for me, when I came from Padan, Rachel died

by me in the land of Canaan in the way, when yet there was but a little way to come unto Ephrath: and I buried

enim quando ueniebam de Mesopotamiam mortua est Rahel her there in the way of Ephrath; the same is Bethlehem. in terra Chanaan in ipso itinere eratque uernum tempus et ingrediebar Ephratam et sepeliui eam iuxta uiam Ephratae quae alio nomine appellatur Bethleem 48:8 uidens autem 48:8 And Israel beheld Joseph’s sons, and said, Who are

these?

filios eius dixit ad eum qui sunt isti 48:9 respondit filii mei 48:9 And Joseph said unto his father, They are my sons,

whom God hath given me in this place. And he said, Bring them, I pray thee, unto me, and I will bless them.

sunt quos dedit mihi Deus in hoc loco adduc inquit eos ad me ut benedicam illis 48:10 oculi enim Israhel caligabant 48:10 Now the eyes of Israel were dim for age, so that he

could not see. And he brought them near unto him; and he kissed them, and embraced them.

prae nimia senectute et clare uidere non poterat adplicitosque ad se deosculatus et circumplexus 48:11 dixit ad filium 48:11 And Israel said unto Joseph, I had not thought to see

thy face: and, lo, God hath shewed me also thy seed.

non sum fraudatus aspectu tuo insuper ostendit mihi Deus semen tuum 48:12 cumque tulisset eos Ioseph de gremio 48:12 And Joseph brought them out from between his

knees, and he bowed himself with his face to the earth.

patris adorauit pronus in terram 48:13 et posuit Ephraim ad 48:13 And Joseph took them both, Ephraim in his right

hand toward Israel’s left hand, and Manasseh in his left hand toward Israel’s right hand, and brought them near

dexteram suam id est ad sinistram Israhel Manassen uero in unto him. sinistra sua ad dexteram scilicet patris adplicuitque ambos

ad eum 48:14 qui extendens manum dextram posuit super 48:14 And Israel stretched out his right hand, and laid it upon Ephraim’s head, who was the younger, and his left hand upon Manasseh’s head, guiding his hands wittingly;

caput Ephraim iunioris fratris sinistram autem super caput for Manasseh was the firstborn.

Manasse qui maior natu erat commutans manus 48:15 be- 48:15 And he blessed Joseph, and said, God, before whom my fathers Abraham and Isaac did walk, the God which fed me all my life long unto this day,

nedixitque Ioseph filio suo et ait Deus in cuius conspectu ambulauerunt patres mei Abraham et Isaac Deus qui pascit me ab adulescentia mea usque in praesentem diem 48:16 48:16 The Angel which redeemed me from all evil, bless

the lads; and let my name be named on them, and the name of my fathers Abraham and Isaac; and let them grow into

angelus qui eruit me de cunctis malis benedicat pueris et a multitude in the midst of the earth. inuocetur super eos nomen meum nomina quoque patrum meorum Abraham et Isaac et crescant in multitudinem super terram 48:17 uidens autem Ioseph quod posuisset pater 48:17 And when Joseph saw that his father laid his right

hand upon the head of Ephraim, it displeased him: and he held up his father’s hand, to remove it from Ephraim’s

suus dexteram manum super caput Ephraim grauiter accepit head unto Manasseh’s head. et adprehensam patris manum leuare conatus est de capite Ephraim et transferre super caput Manasse 48:18 dixitque 48:18 And Joseph said unto his father, Not so, my father:

for this is the firstborn; put thy right hand upon his head.

ad patrem non ita conuenit pater quia hic est primogenitus

pone dexteram tuam super caput eius 48:19 qui rennuens 48:19 And his father refused, and said, I know it, my son, I know it: he also shall become a people, and he also shall be great: but truly his younger brother shall be greater than

ait scio fili mi scio et iste quidem erit in populos et multip- he, and his seed shall become a multitude of nations. licabitur sed frater eius iunior maior illo erit et semen illius

Biblia Sacra Vulgata

48:20 And he blessed them that day, saying, In thee shall crescet in gentes 48:20 benedixitque eis in ipso tempore di-

Israel bless, saying, God make thee as Ephraim and as Manasseh: and he set Ephraim before Manasseh.

cens in te benedicetur Israhel atque dicetur faciat tibi Deus sicut Ephraim et sicut Manasse constituitque Ephraim ante

48:21 And Israel said unto Joseph, Behold, I die: but God Manassen 48:21 et ait ad Ioseph filium suum en ego morior shall be with you, and bring you again unto the land of your fathers.

et erit Deus uobiscum reducetque uos ad terram patrum ue-

48:22 Moreover I have given to thee one portion above thy strorum 48:22 do tibi partem unam extra fratres tuos quam

brethren, which I took out of the hand of the Amorite with my sword and with my bow.

tuli de manu Amorrei in gladio et arcu meo

49:1 And Jacob called unto his sons, and said, Gather 49:1 uocauit autem Iacob filios suos et ait eis congrega-

yourselves together, that I may tell you that which shall befall you in the last days.

mini ut adnuntiem quae uentura sunt uobis diebus nouissi-

49:2 Gather yourselves together, and hear, ye sons of Ja- mis 49:2 congregamini et audite filii Iacob audite Israhel

cob; and hearken unto Israel your father.

49:3 Reuben, thou art my firstborn, my might, and the be- patrem uestrum 49:3 Ruben primogenitus meus tu fortitudo

ginning of my strength, the excellency of dignity, and the excellency of power:

mea et principium doloris mei prior in donis maior impe-

49:4 Unstable as water, thou shalt not excel; because thou rio 49:4 effusus es sicut aqua non crescas quia ascendisti wentest up to thy father’s bed; then defiledst thou it: he went up to my couch.

cubile patris tui et maculasti stratum eius 49:5 Symeon et

49:5 Simeon and Levi are brethren; instruments of cruelty are in their habitations.

49:6 O my soul, come not thou into their secret; unto their Leui fratres uasa iniquitatis bellantia 49:6 in consilio eo-

assembly, mine honour, be not thou united: for in their anger they slew a man, and in their selfwill they digged

down a wall. rum ne ueniat anima mea et in coetu illorum non sit glo-

ria mea quia in furore suo occiderunt uirum et in uoluntate

49:7 Cursed be their anger, for it was fierce; and their sua suffoderunt murum 49:7 maledictus furor eorum quia

wrath, for it was cruel: I will divide them in Jacob, and scatter them in Israel.

pertinax et indignatio illorum quia dura diuidam eos in Ia-

49:8 Judah, thou art he whom thy brethren shall praise: cob et dispergam illos in Israhel 49:8 Iuda te laudabunt fra-

thy hand shall be in the neck of thine enemies; thy father’s children shall bow down before thee.

tres tui manus tua in ceruicibus inimicorum tuorum adora-

49:9 Judah is a lion’s whelp: from the prey, my son, thou bunt te filii patris tui 49:9 catulus leonis Iuda a praeda fili

art gone up: he stooped down, he couched as a lion, and as an old lion; who shall rouse him up?

mi ascendisti requiescens accubuisti ut leo et quasi leaena

49:10 The sceptre shall not depart from Judah, nor a law- quis suscitabit eum 49:10 non auferetur sceptrum de Iuda

giver from between his feet, until Shiloh come; and unto him shall the gathering of the people be.

et dux de femoribus eius donec ueniat qui mittendus est et

49:11 Binding his foal unto the vine, and his ass’s colt unto ipse erit expectatio gentium 49:11 ligans ad uineam pullum the choice vine; he washed his garments in wine, and his clothes in the blood of grapes:

suum et ad uitem o fili mi asinam suam lauabit uino sto-

49:12 His eyes shall be red with wine, and his teeth white lam suam et sanguine uuae pallium suum 49:12 pulchriores

with milk.

49:13 Zebulun shall dwell at the haven of the sea; and he oculi eius uino et dentes lacte candidiores 49:13 Zabulon in

shall be for an haven of ships; and his border shall be unto Zidon.

litore maris habitabit et in statione nauium pertingens usque

49:14 Issachar is a strong ass couching down between two ad Sidonem 49:14 Isachar asinus fortis accubans inter ter-

burdens:

49:15 And he saw that rest was good, and the land that it minos 49:15 uidit requiem quod esset bona et terram quod

was pleasant; and bowed his shoulder to bear, and became a servant unto tribute.

optima et subposuit umerum suum ad portandum factusque

Liber Genesis

est tributis seruiens 49:16 Dan iudicabit populum suum sic- 49:16 Dan shall judge his people, as one of the tribes of

Israel.

ut et alia tribus Israhel 49:17 fiat Dan coluber in uia cerastes 49:17 Dan shall be a serpent by the way, an adder in the

path, that biteth the horse heels, so that his rider shall fall backward.

in semita mordens ungulas equi ut cadat ascensor eius retro 49:18 salutare tuum expectabo Domine 49:19 Gad ac- 49:18 I have waited for thy salvation, O LORD.

49:19 Gad, a troop shall overcome him: but he shall overcome at the last.

cinctus proeliabitur ante eum et ipse accingetur retrorsum 49:20 Aser pinguis panis eius et praebebit delicias regibus 49:20 Out of Asher his bread shall be fat, and he shall

yield royal dainties.

49:21 Nepthalim ceruus emissus et dans eloquia pulchritudi- 49:21 Naphtali is a hind let loose: he giveth goodly words. nis 49:22 filius adcrescens Ioseph filius adcrescens et deco- 49:22 Joseph is a fruitful bough, even a fruitful bough by

a well; whose branches run over the wall:

rus aspectu filiae discurrerunt super murum 49:23 sed ex- 49:23 The archers have sorely grieved him, and shot at

him, and hated him:

asperauerunt eum et iurgati sunt inuideruntque illi habentes iacula 49:24 sedit in forti arcus eius et dissoluta sunt 49:24 But his bow abode in strength, and the arms of his

hands were made strong by the hands of the mighty God of Jacob; (from thence is the shepherd, the stone of Israel:)

uincula brachiorum et manuum illius per manus potentis Iacob inde pastor egressus est lapis Israhel 49:25 Deus patris 49:25 Even by the God of thy father, who shall help thee;

and by the Almighty, who shall bless thee with blessings of heaven above, blessings of the deep that lieth under,

tui erit adiutor tuus et Omnipotens benedicet tibi benedic- blessings of the breasts, and of the womb: tionibus caeli desuper benedictionibus abyssi iacentis deorsum benedictionibus uberum et uuluae 49:26 benedictiones 49:26 The blessings of thy father have prevailed above the

blessings of my progenitors unto the utmost bound of the everlasting hills: they shall be on the head of Joseph, and on the crown of the head of him that was separate from his brethren.

patris tui confortatae sunt benedictionibus patrum eius donec ueniret desiderium collium aeternorum fiant in capite Ioseph et in uertice nazarei inter fratres suos 49:27 Benia- 49:27 Benjamin shall ravin as a wolf: in the morning he

shall devour the prey, and at night he shall divide the spoil.

min lupus rapax mane comedet praedam et uespere diuidet

spolia 49:28 omnes hii in tribubus Israhel duodecim haec 49:28 All these are the twelve tribes of Israel: and this is it that their father spake unto them, and blessed them; every one according to his blessing he blessed them.

locutus est eis pater suus benedixitque singulis benedictionibus propriis 49:29 et praecepit eis dicens ego congregor 49:29 And he charged them, and said unto them, I am to

be gathered unto my people: bury me with my fathers in the cave that is in the field of Ephron the Hittite,

ad populum meum sepelite me cum patribus meis in spelunca duplici quae est in agro Ephron Hetthei 49:30 con- 49:30 In the cave that is in the field of Machpelah, which

is before Mamre, in the land of Canaan, which Abraham bought with the field of Ephron the Hittite for a possession

tra Mambre in terra Chanaan quam emit Abraham cum agro of a buryingplace. ab Ephron Hettheo in possessionem sepulchri 49:31 ibi se- 49:31 There they buried Abraham and Sarah his wife;

there they buried Isaac and Rebekah his wife; and there I buried Leah.

pelierunt eum et Sarram uxorem eius ibi sepultus est Isaac cum Rebecca coniuge ibi et Lia condita iacet 49:32 fini- 49:32 And when Jacob had made an end of commanding

his sons, he gathered up his feet into the bed, and yielded up the ghost, and was gathered unto his people.

tisque mandatis quibus filios instruebat collegit pedes suos super lectulum et obiit adpositusque est ad populum suum

50:1 quod cernens Ioseph ruit super faciem patris flens 50:1 And Joseph fell upon his father’s face, and wept upon

him, and kissed him.

et deosculans eum 50:2 praecepitque seruis suis medicis 50:2 And Joseph commanded his servants the physicians

to embalm his father: and the physicians embalmed Israel.

Biblia Sacra Vulgata

50:3 And forty days were fulfilled for him; for so are ut aromatibus condirent patrem 50:3 quibus iussa explen-

fulfilled the days of those which are embalmed: and the Egyptians mourned for him threescore and ten days.

tibus transierunt quadraginta dies iste quippe mos erat cadauerum conditorum fleuitque eum Aegyptus septuaginta die-

50:4 And when the days of his mourning were past, Joseph bus 50:4 et expleto planctus tempore locutus est Ioseph ad

spake unto the house of Pharaoh, saying, If now I have found grace in your eyes, speak, I pray you, in the ears of

Pharaoh, saying, familiam Pharaonis si inueni gratiam in conspectu uestro lo-50:5 My father made me swear, saying, Lo, I die: in my quimini in auribus Pharaonis 50:5 eo quod pater meus adiu-

grave which I have digged for me in the land of Canaan, there shalt thou bury me. Now therefore let me go up, I

pray thee, and bury my father, and I will come again. rauerit me dicens en morior in sepulchro meo quod fodi mihi

in terra Chanaan sepelies me ascendam igitur et sepeliam pa-

50:6 And Pharaoh said, Go up, and bury thy father, ac- trem meum ac reuertar 50:6 dixitque ei Pharao ascende et

cording as he made thee swear.

50:7 And Joseph went up to bury his father: and with him sepeli patrem tuum sicut adiuratus es 50:7 quo ascendente

went up all the servants of Pharaoh, the elders of his house, and all the elders of the land of Egypt,

ierunt cum eo omnes senes domus Pharaonis cunctique ma-

50:8 And all the house of Joseph, and his brethren, and his iores natu terrae Aegypti 50:8 domus Ioseph cum fratribus

father’s house: only their little ones, and their flocks, and their herds, they left in the land of Goshen.

suis absque paruulis et gregibus atque armentis quae dere-

50:9 And there went up with him both chariots and horse- liquerant in terra Gessen 50:9 habuit quoque in comitatu

men: and it was a very great company.

50:10 And they came to the threshingfloor of Atad, which currus et equites et facta est turba non modica 50:10 ue-

is beyond Jordan, and there they mourned with a great and very sore lamentation: and he made a mourning for his

father seven days. neruntque ad aream Atad quae sita est trans Iordanem ubi

celebrantes exequias planctu magno atque uehementi imp-

50:11 And when the inhabitants of the land, the Canaan- leuerunt septem dies 50:11 quod cum uidissent habitatores

ites, saw the mourning in the floor of Atad, they said, This is a grievous mourning to the Egyptians: wherefore the name of it was called Abelmizraim, which is beyond Jordan.

terrae Chanaan dixerunt planctus magnus est iste Aegyptiis et idcirco appellauerunt nomen loci illius Planctus Aegypti

50:12 And his sons did unto him according as he com- 50:12 fecerunt ergo filii Iacob sicut praeceperat eis 50:13 et

manded them: 50:13 For his sons carried him into the land of Canaan, and buried him in the cave of the field of Machpelah, which Abraham bought with the field for a possession of a buryingplace of Ephron the Hittite, before Mamre.

portantes eum in terram Chanaan sepelierunt in spelunca duplici quam emerat Abraham cum agro in possessionem se-

50:14 And Joseph returned into Egypt, he, and his pulchri ab Ephron Hettheo contra faciem Mambre 50:14 re-

brethren, and all that went up with him to bury his father, after he had buried his father.

uersusque est Ioseph in Aegyptum cum fratribus suis et omni

50:15 And when Joseph’s brethren saw that their father comitatu sepulto patre 50:15 quo mortuo timentes fratres

was dead, they said, Joseph will peradventure hate us, and will certainly requite us all the evil which we did unto him.

eius et mutuo conloquentes ne forte memor sit iniuriae quam

50:16 And they sent a messenger unto Joseph, saying, Thy passus est et reddat nobis malum omne quod fecimus 50:16 father did command before he died, saying,

mandauerunt ei pater tuus praecepit nobis antequam more-

50:17 So shall ye say unto Joseph, Forgive, I pray thee retur 50:17 ut haec tibi uerbis illius diceremus obsecro ut

now, the trespass of thy brethren, and their sin; for they did unto thee evil: and now, we pray thee, forgive the trespass of the servants of the God of thy father. And Joseph wept when they spake unto him.

obliuiscaris sceleris fratrum tuorum et peccati atque malitiae quam exercuerunt in te nos quoque oramus ut seruis Dei patris tui dimittas iniquitatem hanc quibus auditis fleuit Io-

Liber Genesis

seph 50:18 ueneruntque ad eum fratres sui et proni in ter- 50:18 And his brethren also went and fell down before his

face; and they said, Behold, we be thy servants.

ram dixerunt serui tui sumus 50:19 quibus ille respondit no- 50:19 And Joseph said unto them, Fear not: for am I in the

place of God?

lite timere num Dei possumus rennuere uoluntatem 50:20 50:20 But as for you, ye thought evil against me; but God

meant it unto good, to bring to pass, as it is this day, to save much people alive.

uos cogitastis de me malum et Deus uertit illud in bonum ut exaltaret me sicut inpraesentiarum cernitis et saluos faceret multos populos 50:21 nolite metuere ego pascam uos et 50:21 Now therefore fear ye not: I will nourish you, and

your little ones. And he comforted them, and spake kindly unto them.

paruulos uestros consolatusque est eos et blande ac leniter est locutus 50:22 et habitauit in Aegypto cum omni domo 50:22 And Joseph dwelt in Egypt, he, and his father’s

house: and Joseph lived an hundred and ten years; and Joseph saw Ephraim’s children of the third generation: the children also of Machir the son of Manasseh were brought up upon Joseph’s knees.

patris sui uixitque centum decem annis et uidit Ephraim filios usque ad tertiam generationem filii quoque Machir filii Manasse nati sunt in genibus Ioseph 50:23 quibus transactis 50:23 And Joseph said unto his brethren, I die: and God

will surely visit you, and bring you out of this land unto the land which he sware to Abraham, to Isaac, and to Jacob.

locutus est fratribus suis post mortem meam Deus uisitabit uos et ascendere faciet de terra ista ad terram quam iurauit Abraham Isaac et Iacob 50:24 cumque adiurasset eos at- 50:24 And Joseph took an oath of the children of Israel,

saying, God will surely visit you, and ye shall carry up my bones from hence.

que dixisset Deus uisitabit uos asportate uobiscum ossa mea de loco isto 50:25 mortuus est expletis centum decem uitae 50:25 So Joseph died, being an hundred and ten years old:

and they embalmed him, and he was put in a coffin in Egypt.

suae annis et conditus aromatibus repositus est in loculo in Aegypto

Biblia Sacra Vulgata

XC

\underline{Liber Exodus}

1:1 haec sunt nomina filiorum Israhel qui ingressi sunt 1:1 Now these are the names of the children of Israel,

which came into Egypt; every man and his household came with Jacob.

Aegyptum cum Iacob singuli cum domibus suis introierunt 1:2 Ruben Symeon Leui Iuda 1:3 Isachar Zabulon et Be- 1:2 Reuben, Simeon, Levi, and Judah,

1:3 Issachar, Zebulun, and Benjamin,

niamin 1:4 Dan et Nepthalim Gad et Aser 1:5 erant igi- _{1:4 Dan, and Naphtali, Gad, and Asher.}

1:5 And all the souls that came out of the loins of Jacob were seventy souls: for Joseph was in Egypt already.

tur omnes animae eorum qui egressi sunt de femore Iacob

septuaginta Ioseph autem in Aegypto erat 1:6 quo mortuo 1:6 And Joseph died, and all his brethren, and all that gen-

eration.

et uniuersis fratribus eius omnique cognatione illa 1:7 filii 1:7 And the children of Israel were fruitful, and increased

abundantly, and multiplied, and waxed exceeding mighty; and the land was filled with them.

Israhel creuerunt et quasi germinantes multiplicati sunt ac roborati nimis impleuerunt terram 1:8 surrexit interea rex 1:8 Now there arose up a new king over Egypt, which

knew not Joseph.

nouus super Aegyptum qui ignorabat Ioseph 1:9 et ait ad 1:9 And he said unto his people, Behold, the people of the

children of Israel are more and mightier than we:

populum suum ecce populus filiorum Israhel multus et fortior nobis 1:10 uenite sapienter opprimamus eum ne forte 1:10 Come on, let us deal wisely with them; lest they mul-

tiply, and it come to pass, that, when there falleth out any war, they join also unto our enemies, and fight against us,

multiplicetur et si ingruerit contra nos bellum addatur in- and so get them up out of the land. imicis nostris expugnatisque nobis egrediatur e terra 1:11 1:11 Therefore they did set over them taskmasters to af-

flict them with their burdens. And they built for Pharaoh treasure cities, Pithom and Raamses.

praeposuit itaque eis magistros operum ut adfligerent eos oneribus aedificaueruntque urbes tabernaculorum Pharaoni Phiton et Ramesses 1:12 quantoque opprimebant eos tanto 1:12 But the more they afflicted them, the more they mul-

tiplied and grew. And they were grieved because of the children of Israel.

magis multiplicabantur et crescebant 1:13 oderantque filios

1:13 And the Egyptians made the children of Israel to serve with rigour:

Israhel Aegyptii et adfligebant inludentes eis 1:14 atque ad 1:14 And they made their lives bitter with hard bondage,

in morter, and in brick, and in all manner of service in the field: all their service, wherein they made them serve, was

amaritudinem perducebant uitam eorum operibus duris luti with rigour. et lateris omnique famulatu quo in terrae operibus premebantur 1:15 dixit autem rex Aegypti obsetricibus Hebraeo- 1:15 And the king of Egypt spake to the Hebrew mid-

wives, of which the name of the one was Shiphrah, and the name of the other Puah:

rum quarum una uocabatur Sephra altera Phua 1:16 prae-

1:16 And he said, When ye do the office of a midwife to the Hebrew women, and see them upon the stools; if it be a son, then ye shall kill him: but if it be a daughter, then she shall live.

cipiens eis quando obsetricabitis Hebraeas et partus tempus aduenerit si masculus fuerit interficite illum si femina reseruate 1:17 timuerunt autem obsetrices Deum et non fece- 1:17 But the midwives feared God, and did not as the king

of Egypt commanded them, but saved the men children alive.

runt iuxta praeceptum regis Aegypti sed conseruabant mares 1:18 quibus ad se accersitis rex ait quidnam est hoc quod fa- 1:18 And the king of Egypt called for the midwives, and

said unto them, Why have ye done this thing, and have saved the men children alive?

cere uoluistis ut pueros seruaretis 1:19 quae responderunt

1:19 And the midwives said unto Pharaoh, Because the Hebrew women are not as the Egyptian women; for they are lively, and are delivered ere the midwives come in unto them.

non sunt hebraeae sicut aegyptiae mulieres ipsae enim obsetricandi habent scientiam et priusquam ueniamus ad eas pariunt 1:20 bene ergo fecit Deus obsetricibus et creuit po- 1:20 Therefore God dealt well with the midwives: and the

people multiplied, and waxed very mighty.

pulus confortatusque est nimis 1:21 et quia timuerant ob- 1:21 And it came to pass, because the midwives feared

God, that he made them houses.

Biblia Sacra Vulgata

1:22 And Pharaoh charged all his people, saying, Every setrices Deum aedificauit illis domos 1:22 praecepit autem

son that is born ye shall cast into the river, and every daughter ye shall save alive.

Pharao omni populo suo dicens quicquid masculini sexus natum fuerit in flumen proicite quicquid feminei reseruate

2:1 And there went a man of the house of Levi, and took 2:1 egressus est post haec uir de domo Leui accepta ux-

to wife a daughter of Levi.

2:2 And the woman conceived, and bare a son: and when ore stirpis suae 2:2 quae concepit et peperit filium et ui-

she saw him that he was a goodly child, she hid him three months.

dens eum elegantem abscondit tribus mensibus 2:3 cumque

2:3 And when she could not longer hide him, she took for him an ark of bulrushes, and daubed it with slime and with pitch, and put the child therein; and she laid it in the flags by the river’s brink.

iam celare non posset sumpsit fiscellam scirpeam et liniuit eam bitumine ac pice posuitque intus infantulum et expo-

2:4 And his sister stood afar off, to wit what would be done suit eum in carecto ripae fluminis 2:4 stante procul sorore

to him.

2:5 And the daughter of Pharaoh came down to wash her- eius et considerante euentum rei 2:5 ecce autem descen-

self at the river; and her maidens walked along by the river’s side; and when she saw the ark among the flags,

she sent her maid to fetch it. debat filia Pharaonis ut lauaretur in flumine et puellae eius

gradiebantur per crepidinem aluei quae cum uidisset fiscel-

2:6 And when she had opened it, she saw the child: and, lam in papyrione misit unam e famulis suis et adlatam 2:6

behold, the babe wept. And she had compassion on him, and said, This is one of the Hebrews’ children.

aperiens cernensque in ea paruulum uagientem miserta eius

2:7 Then said his sister to Pharaoh’s daughter, Shall I go ait de infantibus Hebraeorum est 2:7 cui soror pueri uis in-

and call to thee a nurse of the Hebrew women, that she may nurse the child for thee?

quit ut uadam et uocem tibi hebraeam mulierem quae nut-

2:8 And Pharaoh’s daughter said to her, Go. And the maid rire possit infantulum 2:8 respondit uade perrexit puella et

went and called the child’s mother. 2:9 And Pharaoh’s daughter said unto her, Take this child uocauit matrem eius 2:9 ad quam locuta filia Pharaonis acaway, and nurse it for me, and I will give thee thy wages. And the women took the child, and nursed it.

cipe ait puerum istum et nutri mihi ego tibi dabo mercedem tuam suscepit mulier et nutriuit puerum adultumque tradi-

2:10 And the child grew, and she brought him unto dit filiae Pharaonis 2:10 quem illa adoptauit in locum filii

Pharaoh’s daughter, and he became her son. And she called his name Moses: and she said, Because I drew him

out of the water. uocauitque nomen eius Mosi dicens quia de aqua tuli eum 2:11 And it came to pass in those days, when Moses was 2:11 in diebus illis postquam creuerat Moses egressus ad fra-

grown, that he went out unto his brethren, and looked on their burdens: and he spied an Egyptian smiting an He-

brew, one of his brethren. tres suos uidit adflictionem eorum et uirum aegyptium per-2:12 And he looked this way and that way, and when he cutientem quendam de Hebraeis fratribus suis 2:12 cum-

saw that there was no man, he slew the Egyptian, and hid him in the sand.

que circumspexisset huc atque illuc et nullum adesse uidis-

2:13 And when he went out the second day, behold, two set percussum Aegyptium abscondit sabulo 2:13 et egres-

men of the Hebrews strove together: and he said to him that did the wrong, Wherefore smitest thou thy fellow?

sus die altero conspexit duos Hebraeos rixantes dixitque ei

2:14 And he said, Who made thee a prince and a judge qui faciebat iniuriam quare percutis proximum tuum 2:14

over us? intendest thou to kill me, as thou killedst the Egyptian? And Moses feared, and said, Surely this thing

is known. qui respondit quis constituit te principem et iudicem super nos num occidere me tu dicis sicut occidisti Aegyptium timuit Moses et ait quomodo palam factum est uerbum istud

2:15 Now when Pharaoh heard this thing, he sought to 2:15 audiuitque Pharao sermonem hunc et quaerebat occi-

slay Moses. But Moses fled from the face of Pharaoh, and dwelt in the land of Midian: and he sat down by a well.

dere Mosen qui fugiens de conspectu eius moratus est in terra Madian et sedit iuxta puteum 2:16 erant sacerdoti Ma- 2:16 Now the priest of Midian had seven daughters: and

they came and drew water, and filled the troughs to water their father’s flock.

dian septem filiae quae uenerunt ad hauriendas aquas et impletis canalibus adaquare cupiebant greges patris sui 2:17 2:17 And the shepherds came and drove them away: but

Moses stood up and helped them, and watered their flock.

superuenere pastores et eiecerunt eas surrexitque Moses et defensis puellis adaquauit oues earum 2:18 quae cum re- 2:18 And when they came to Reuel their father, he said,

How is it that ye are come so soon to day?

uertissent ad Raguhel patrem suum dixit ad eas cur uelocius uenistis solito 2:19 responderunt uir aegyptius liberauit nos 2:19 And they said, An Egyptian delivered us out of the

hand of the shepherds, and also drew water enough for us, and watered the flock.

de manu pastorum insuper et hausit aquam nobiscum potumque dedit ouibus 2:20 at ille ubi est inquit quare dimi- 2:20 And he said unto his daughters, And where is he?

why is it that ye have left the man? call him, that he may eat bread.

sistis hominem uocate eum ut comedat panem 2:21 iurauit

2:21 And Moses was content to dwell with the man: and he gave Moses Zipporah his daughter.

ergo Moses quod habitaret cum eo accepitque Sefforam filiam eius 2:22 quae peperit filium quem uocauit Gersam di- 2:22 And she bare him a son, and he called his name Ger-

shom: for he said, I have been a stranger in a strange land.

cens aduena fui in terra aliena 2:23 post multum temporis 2:23 And it came to pass in process of time, that the king

of Egypt died: and the children of Israel sighed by reason of the bondage, and they cried, and their cry came up unto

mortuus est rex Aegypti et ingemescentes filii Israhel propter God by reason of the bondage. opera uociferati sunt ascenditque clamor eorum ad Deum ab operibus 2:24 et audiuit gemitum eorum ac recordatus foe- 2:24 And God heard their groaning, and God remembered

his covenant with Abraham, with Isaac, and with Jacob.

deris quod pepigerat cum Abraham et Isaac et Iacob 2:25 2:25 And God looked upon the children of Israel, and God

had respect unto them.

respexit filios Israhel et cognouit eos

3:1 Moses autem pascebat oues Iethro cognati sui sacer- 3:1 Now Moses kept the flock of Jethro his father in law,

the priest of Midian: and he led the flock to the backside of the desert, and came to the mountain of God, even to

dotis Madian cumque minasset gregem ad interiora deserti Horeb. uenit ad montem Dei Horeb 3:2 apparuitque ei Dominus 3:2 And the angel of the LORD appeared unto him in a

flame of fire out of the midst of a bush: and he looked, and, behold, the bush burned with fire, and the bush was

in flamma ignis de medio rubi et uidebat quod rubus arde- not consumed. ret et non conbureretur 3:3 dixit ergo Moses uadam et ui- 3:3 And Moses said, I will now turn aside, and see this

great sight, why the bush is not burnt.

debo uisionem hanc magnam quare non conburatur rubus 3:4 cernens autem Dominus quod pergeret ad uidendum uo- 3:4 And when the LORD saw that he turned aside to see,

God called unto him out of the midst of the bush, and said, Moses, Moses. And he said, Here am I.

cauit eum de medio rubi et ait Moses Moses qui respondit adsum 3:5 at ille ne adpropies inquit huc solue calciamen- 3:5 And he said, Draw not nigh hither: put off thy shoes

from off thy feet, for the place whereon thou standest is holy ground.

tum de pedibus tuis locus enim in quo stas terra sancta est 3:6 et ait ego sum Deus patris tui Deus Abraham Deus Isaac 3:6 Moreover he said, I am the God of thy father, the God

of Abraham, the God of Isaac, and the God of Jacob. And Moses hid his face; for he was afraid to look upon God.

Deus Iacob abscondit Moses faciem suam non enim audebat aspicere contra Deum 3:7 cui ait Dominus uidi adflictio- 3:7 And the LORD said, I have surely seen the affliction

of my people which are in Egypt, and have heard their cry by reason of their taskmasters; for I know their sorrows;

nem populi mei in Aegypto et clamorem eius audiui propter 3:8 And I am come down to deliver them out of the hand of duritiam eorum qui praesunt operibus 3:8 et sciens dolo-

the Egyptians, and to bring them up out of that land unto a good land and a large, unto a land flowing with milk and honey; unto the place of the Canaanites, and the Hittites, and the Amorites, and the Perizzites, and the Hivites, and the Jebusites.

rem eius descendi ut liberarem eum de manibus Aegyptiorum et educerem de terra illa in terram bonam et spatiosam in terram quae fluit lacte et melle ad loca Chananei et Het-

3:9 Now therefore, behold, the cry of the children of Is- thei et Amorrei Ferezei et Euei et Iebusei 3:9 clamor ergo

rael is come unto me: and I have also seen the oppression wherewith the Egyptians oppress them.

filiorum Israhel uenit ad me uidique adflictionem eorum qua

3:10 Come now therefore, and I will send thee unto ab Aegyptiis opprimuntur 3:10 sed ueni mittam te ad Pha-

Pharaoh, that thou mayest bring forth my people the children of Israel out of Egypt.

raonem ut educas populum meum filios Israhel de Aegypto

3:11 And Moses said unto God, Who am I, that I should 3:11 dixit Moses ad Deum quis ego sum ut uadam ad Pha-

go unto Pharaoh, and that I should bring forth the children of Israel out of Egypt?

raonem et educam filios Israhel de Aegypto 3:12 qui dixit

3:12 And he said, Certainly I will be with thee; and this shall be a token unto thee, that I have sent thee: When thou hast brought forth the people out of Egypt, ye shall serve God upon this mountain.

ei ero tecum et hoc habebis signum quod miserim te cum eduxeris populum de Aegypto immolabis Deo super mon-

3:13 And Moses said unto God, Behold, when I come unto tem istum 3:13 ait Moses ad Deum ecce ego uadam ad fithe children of Israel, and shall say unto them, The God of your fathers hath sent me unto you; and they shall say to

me, What is his name? what shall I say unto them? lios Israhel et dicam eis Deus patrum uestrorum misit me ad uos si dixerint mihi quod est nomen eius quid dicam eis

3:14 And God said unto Moses, I AM THAT I AM: and 3:14 dixit Deus ad Mosen ego sum qui sum ait sic dices filiis

he said, Thus shalt thou say unto the children of Israel, I AM hath sent me unto you.

Israhel qui est misit me ad uos 3:15 dixitque iterum Deus

3:15 And God said moreover unto Moses, Thus shalt thou say unto the children of Israel, the LORD God of your fathers, the God of Abraham, the God of Isaac, and the God of Jacob, hath sent me unto you: this is my name for ever, and this is my memorial unto all generations.

ad Mosen haec dices filiis Israhel Dominus Deus patrum uestrorum Deus Abraham Deus Isaac et Deus Iacob misit me ad uos hoc nomen mihi est in aeternum et hoc memoriale

3:16 Go, and gather the elders of Israel together, and say meum in generationem et generatione 3:16 uade congrega unto them, The LORD God of your fathers, the God of Abraham, of Isaac, and of Jacob, appeared unto me, saying, I have surely visited you, and seen that which is done to you in Egypt:

seniores Israhel et dices ad eos Dominus Deus patrum uestrorum apparuit mihi Deus Abraham et Deus Isaac et Deus Iacob dicens uisitans uisitaui uos et omnia quae acciderunt

3:17 And I have said, I will bring you up out of the afflic- uobis in Aegypto 3:17 et dixi ut educam uos de adflictione

tion of Egypt unto the land of the Canaanites, and the Hittites, and the Amorites, and the Perizzites, and the Hivites, and the Jebusites, unto a land flowing with milk and honey. Aegypti in terram Chananei et Hetthei et Amorrei Ferezei

3:18 And they shall hearken to thy voice: and thou shalt et Euei et Iebusei ad terram fluentem lacte et melle 3:18 et

come, thou and the elders of Israel, unto the king of Egypt, and ye shall say unto him, The LORD God of the Hebrews hath met with us: and now let us go, we beseech thee, three days’ journey into the wilderness, that we may sacrifice to the LORD our God.

audient uocem tuam ingredierisque tu et seniores Israhel ad regem Aegypti et dices ad eum Dominus Deus Hebraeorum uocauit nos ibimus uiam trium dierum per solitudinem ut im-

3:19 And I am sure that the king of Egypt will not let you molemus Domino Deo nostro 3:19 sed ego scio quod non go, no, not by a mighty hand.

dimittet uos rex Aegypti ut eatis nisi per manum ualidam

3:20 And I will stretch out my hand, and smite Egypt with 3:20 extendam enim manum meam et percutiam Aegyptum all my wonders which I will do in the midst thereof: and after that he will let you go.

in cunctis mirabilibus meis quae facturus sum in medio eo-

rum post haec dimittet uos 3:21 daboque gratiam populo 3:21 And I will give this people favour in the sight of the Egyptians: and it shall come to pass, that, when ye go, ye shall not go empty.

huic coram Aegyptiis et cum egrediemini non exibitis uacui 3:22 sed postulabit mulier a uicina sua et ab hospita uasa ar- 3:22 But every woman shall borrow of her neighbour, and

of her that sojourneth in her house, jewels of silver, and jewels of gold, and raiment: and ye shall put them upon your sons, and upon your daughters; and ye shall spoil the Egyptians.

gentea et aurea ac uestes ponetisque eas super filios et filias uestras et spoliabitis Aegyptum

4:1 respondens Moses ait non credent mihi neque audient 4:1 And Moses answered and said, But, behold, they will

not believe me, nor hearken unto my voice: for they will say, The LORD hath not appeared unto thee.

uocem meam sed dicent non apparuit tibi Dominus 4:2

4:2 And the LORD said unto him, What is that in thine hand? And he said, A rod.

dixit ergo ad eum quid est hoc quod tenes in manu tua respondit uirga 4:3 ait proice eam in terram proiecit et uersa 4:3 And he said, Cast it on the ground. And he cast it on

the ground, and it became a serpent; and Moses fled from before it.

est in colubrum ita ut fugeret Moses 4:4 dixitque Dominus

4:4 And the LORD said unto Moses, Put forth thine hand, and take it by the tail. And he put forth his hand, and caught it, and it became a rod in his hand:

extende manum tuam et adprehende caudam eius extendit et tenuit uersaque est in uirgam 4:5 ut credant inquit quod ap- 4:5 That they may believe that the LORD God of their

fathers, the God of Abraham, the God of Isaac, and the God of Jacob, hath appeared unto thee.

paruerit tibi Dominus Deus patrum tuorum Deus Abraham Deus Isaac Deus Iacob 4:6 dixitque Dominus rursum mitte 4:6 And the LORD said furthermore unto him, Put now

thine hand into thy bosom. And he put his hand into his bosom: and when he took it out, behold, his hand was

manum in sinum tuum quam cum misisset in sinum pro- leprous as snow. tulit leprosam instar niuis 4:7 retrahe ait manum in sinum 4:7 And he said, Put thine hand into thy bosom again. And

he put his hand into his bosom again; and plucked it out of his bosom, and, behold, it was turned again as his other

tuum retraxit et protulit iterum et erat similis carni reliquae flesh. 4:8 si non crediderint inquit tibi neque audierint sermonem 4:8 And it shall come to pass, if they will not believe thee,

neither hearken to the voice of the first sign, that they will believe the voice of the latter sign.

signi prioris credent uerbo signi sequentis 4:9 quod si nec

4:9 And it shall come to pass, if they will not believe also these two signs, neither hearken unto thy voice, that thou shalt take of the water of the river, and pour it upon the dry land: and the water which thou takest out of the river shall become blood upon the dry land.

duobus quidem his signis crediderint neque audierint uocem tuam sume aquam fluminis et effunde eam super aridam et quicquid hauseris de fluuio uertetur in sanguinem 4:10 ait 4:10 And Moses said unto the LORD, O my LORD, I am

not eloquent, neither heretofore, nor since thou hast spoken unto thy servant: but I am slow of speech, and of a

Moses obsecro Domine non sum eloquens ab heri et nudius slow tongue. tertius et ex quo locutus es ad seruum tuum inpeditioris et tardioris linguae sum 4:11 dixit Dominus ad eum quis fe- 4:11 And the LORD said unto him, Who hath made man’s

mouth? or who maketh the dumb, or deaf, or the seeing, or the blind? have not I the LORD?

cit os hominis aut quis fabricatus est mutum et surdum uidentem et caecum nonne ego 4:12 perge igitur et ego ero 4:12 Now therefore go, and I will be with thy mouth, and

teach thee what thou shalt say.

in ore tuo doceboque te quid loquaris 4:13 at ille obsecro 4:13 And he said, O my LORD, send, I pray thee, by the

hand of him whom thou wilt send.

inquit Domine mitte quem missurus es 4:14 iratus Domi- 4:14 And the anger of the LORD was kindled against

Moses, and he said, Is not Aaron the Levite thy brother? I know that he can speak well. And also, behold, he cometh forth to meet thee: and when he seeth thee, he will be glad in his heart.

nus in Mosen ait Aaron frater tuus Leuites scio quod eloquens sit ecce ipse egreditur in occursum tuum uidensque te laetabitur corde 4:15 loquere ad eum et pone uerba mea in 4:15 And thou shalt speak unto him, and put words in his

mouth: and I will be with thy mouth, and with his mouth, and will teach you what ye shall do.

Biblia Sacra Vulgata

ore eius ego ero in ore tuo et in ore illius et ostendam uobis

4:16 And he shall be thy spokesman unto the people: and quid agere debeatis 4:16 ipse loquetur pro te ad populum et

he shall be, even he shall be to thee instead of a mouth, and thou shalt be to him instead of God.

erit os tuum tu autem eris ei in his quae ad Deum pertinent

4:17 And thou shalt take this rod in thine hand, wherewith 4:17 uirgam quoque hanc sume in manu tua in qua facturus thou shalt do signs.

4:18 And Moses went and returned to Jethro his father in es signa 4:18 abiit Moses et reuersus est ad Iethro cogna-

law, and said unto him, Let me go, I pray thee, and return unto my brethren which are in Egypt, and see whether they

be yet alive. And Jethro said to Moses, Go in peace. tum suum dixitque ei uadam et reuertar ad fratres meos in

Aegyptum ut uideam si adhuc uiuunt cui ait Iethro uade in

4:19 And the LORD said unto Moses in Midian, Go, re- pace 4:19 dixit ergo Dominus ad Mosen in Madian uade re-

turn into Egypt: for all the men are dead which sought thy life.

uertere in Aegyptum mortui sunt omnes qui quaerebant ani-

4:20 And Moses took his wife and his sons, and set them mam tuam 4:20 tulit Moses uxorem et filios suos et inposuit

upon an ass, and he returned to the land of Egypt: and Moses took the rod of God in his hand.

eos super asinum reuersusque est in Aegyptum portans uir-

4:21 And the LORD said unto Moses, When thou goest gam Dei in manu sua 4:21 dixitque ei Dominus reuertenti

to return into Egypt, see that thou do all those wonders before Pharaoh, which I have put in thine hand: but I will

harden his heart, that he shall not let the people go. in Aegyptum uide ut omnia ostenta quae posui in manu tua

facias coram Pharaone ego indurabo cor eius et non dimittet

4:22 And thou shalt say unto Pharaoh, Thus saith the populum 4:22 dicesque ad eum haec dicit Dominus filius

LORD, Israel is my son, even my firstborn:

4:23 And I say unto thee, Let my son go, that he may serve meus primogenitus meus Israhel 4:23 dixi tibi dimitte fi-

me: and if thou refuse to let him go, behold, I will slay thy son, even thy firstborn.

lium meum ut seruiat mihi et noluisti dimittere eum ecce ego

4:24 And it came to pass by the way in the inn, that the interficiam filium tuum primogenitum 4:24 cumque esset in

LORD met him, and sought to kill him.

itinere in diuersorio occurrit ei Dominus et uolebat occidere

4:25 Then Zipporah took a sharp stone, and cut off the eum 4:25 tulit ilico Seffora acutissimam petram et circum-

foreskin of her son, and cast it at his feet, and said, Surely a bloody husband art thou to me.

cidit praeputium filii sui tetigitque pedes eius et ait sponsus

4:26 So he let him go: then she said, A bloody husband sanguinum tu mihi es 4:26 et dimisit eum postquam dixe-

thou art, because of the circumcision.

4:27 And the LORD said to Aaron, Go into the wilderness rat sponsus sanguinum ob circumcisionem 4:27 dixit autem

to meet Moses. And he went, and met him in the mount of God, and kissed him.

Dominus ad Aaron uade in occursum Mosi in deserto qui

4:28 And Moses told Aaron all the words of the LORD perrexit ei obuiam in montem Dei et osculatus est eum 4:28

who had sent him, and all the signs which he had commanded him.

narrauitque Moses Aaron omnia uerba Domini quibus mise-

4:29 And Moses and Aaron went and gathered together all rat eum et signa quae mandauerat 4:29 ueneruntque simul

the elders of the children of Israel:

4:30 And Aaron spake all the words which the LORD had et congregauerunt cunctos seniores filiorum Israhel 4:30 lo-

spoken unto Moses, and did the signs in the sight of the people.

cutusque est Aaron omnia uerba quae dixerat Dominus ad

4:31 And the people believed: and when they heard that Mosen et fecit signa coram populo 4:31 et credidit populus

the LORD had visited the children of Israel, and that he had looked upon their affliction, then they bowed their

heads and worshipped. audieruntque quod uisitasset Dominus filios Israhel et quod respexisset adflictionem eorum et proni adorauerunt

5:1 And afterward Moses and Aaron went in, and told 5:1 post haec ingressi sunt Moses et Aaron et dixerunt

Pharaoh, Thus saith the LORD God of Israel, Let my people go, that they may hold a feast unto me in the wilderness.

Pharaoni haec dicit Dominus Deus Israhel dimitte populum meum ut sacrificet mihi in deserto 5:2 at ille respondit quis 5:2 And Pharaoh said, Who is the LORD, that I should

obey his voice to let Israel go? I know not the LORD, neither will I let Israel go.

est Dominus ut audiam uocem eius et dimittam Israhel nescio Dominum et Israhel non dimittam 5:3 dixerunt Deus 5:3 And they said, The God of the Hebrews hath met with

us: let us go, we pray thee, three days’ journey into the desert, and sacrifice unto the LORD our God; lest he fall

Hebraeorum uocauit nos ut eamus uiam trium dierum in so- upon us with pestilence, or with the sword. litudinem et sacrificemus Domino Deo nostro ne forte accidat nobis pestis aut gladius 5:4 ait ad eos rex Aegypti quare 5:4 And the king of Egypt said unto them, Wherefore do

ye, Moses and Aaron, let the people from their works? get you unto your burdens.

Moses et Aaron sollicitatis populum ab operibus suis ite ad onera uestra 5:5 dixitque Pharao multus est populus terrae 5:5 And Pharaoh said, Behold, the people of the land now

are many, and ye make them rest from their burdens.

uidetis quod turba succreuerit quanto magis si dederitis eis requiem ab operibus 5:6 praecepit ergo in die illo praefectis 5:6 And Pharaoh commanded the same day the taskmas-

ters of the people, and their officers, saying,

operum et exactoribus populi dicens 5:7 nequaquam ultra 5:7 Ye shall no more give the people straw to make brick,

as heretofore: let them go and gather straw for themselves.

dabitis paleas populo ad conficiendos lateres sicut prius sed ipsi uadant et colligant stipulam 5:8 et mensuram laterum 5:8 And the tale of the bricks, which they did make hereto-

fore, ye shall lay upon them; ye shall not diminish ought thereof: for they be idle; therefore they cry, saying, Let us

quos prius faciebant inponetis super eos nec minuetis quic- go and sacrifice to our God. quam uacant enim et idcirco uociferantur dicentes eamus et sacrificemus Deo nostro 5:9 opprimantur operibus et exple- 5:9 Let there more work be laid upon the men, that they

may labour therein; and let them not regard vain words.

ant ea ut non adquiescant uerbis mendacibus 5:10 igitur eg- 5:10 And the taskmasters of the people went out, and their

officers, and they spake to the people, saying, Thus saith Pharaoh, I will not give you straw.

ressi praefecti operum et exactores ad populum dixerunt sic dicit Pharao non do uobis paleas 5:11 ite et colligite sicubi 5:11 Go ye, get you straw where ye can find it: yet not

ought of your work shall be diminished.

inuenire potueritis nec minuetur quicquam de opere uestro 5:12 dispersusque est populus per omnem terram Aegypti ad 5:12 So the people were scattered abroad throughout all

the land of Egypt to gather stubble instead of straw.

colligendas paleas 5:13 praefecti quoque operum instabant 5:13 And the taskmasters hasted them, saying, Fulfil your

works, your daily tasks, as when there was straw.

dicentes conplete opus uestrum cotidie ut prius facere solebatis quando dabantur uobis paleae 5:14 flagellatique sunt 5:14 And the officers of the children of Israel, which

Pharaoh’s taskmasters had set over them, were beaten, and demanded, Wherefore have ye not fulfilled your task in

qui praeerant operibus filiorum Israhel ab exactoribus Pha- making brick both yesterday and to day, as heretofore? raonis dicentibus quare non impletis mensuram laterum sicut prius nec heri nec hodie 5:15 ueneruntque praepositi fi- 5:15 Then the officers of the children of Israel came and

cried unto Pharaoh, saying, Wherefore dealest thou thus with thy servants?

liorum Israhel et uociferati sunt ad Pharaonem dicentes cur ita agis contra seruos tuos 5:16 paleae non dantur nobis et 5:16 There is no straw given unto thy servants, and they

say to us, Make brick: and, behold, thy servants are beaten; but the fault is in thine own people.

lateres similiter imperantur en famuli tui flagellis caedimur et iniuste agitur contra populum tuum 5:17 qui ait uacatis 5:17 But he said, Ye are idle, ye are idle: therefore ye say,

Let us go and do sacrifice to the LORD.

otio et idcirco dicitis eamus et sacrificemus Domino 5:18 5:18 Go therefore now, and work; for there shall no straw

be given you, yet shall ye deliver the tale of bricks.

Biblia Sacra Vulgata

ite ergo et operamini paleae non dabuntur uobis et reddetis

5:19 And the officers of the children of Israel did see that consuetum numerum laterum 5:19 uidebantque se praepothey were in evil case, after it was said, Ye shall not minish ought from your bricks of your daily task.

siti filiorum Israhel in malo eo quod diceretur eis non mi-

5:20 And they met Moses and Aaron, who stood in the nuetur quicquam de lateribus per singulos dies 5:20 occur-

way, as they came forth from Pharaoh:

reruntque Mosi et Aaron qui stabant ex aduerso egredientes

5:21 And they said unto them, The LORD look upon you, a Pharaone 5:21 et dixerunt ad eos uideat Dominus et iudi-

and judge; because ye have made our savour to be abhorred in the eyes of Pharaoh, and in the eyes of his ser-

vants, to put a sword in their hand to slay us. cet quoniam fetere fecistis odorem nostrum coram Pharao et 5:22 And Moses returned unto the LORD, and said, seruis eius et praebuistis ei gladium ut occideret nos 5:22

LORD, wherefore hast thou so evil entreated this people? why is it that thou hast sent me?

reuersusque Moses ad Dominum ait Domine cur adflixisti

5:23 For since I came to Pharaoh to speak in thy name, he populum istum quare misisti me 5:23 ex eo enim quo inhath done evil to this people; neither hast thou delivered thy people at all.

gressus sum ad Pharaonem ut loquerer nomine tuo adflixit populum tuum et non liberasti eos

6:1 Then the LORD said unto Moses, Now shalt thou see 6:1 dixit Dominus ad Mosen nunc uidebis quae facturus

what I will do to Pharaoh: for with a strong hand shall he let them go, and with a strong hand shall he drive them out

of his land. sum Pharaoni per manum enim fortem dimittet eos et in 6:2 And God spake unto Moses, and said unto him, I am manu robusta eiciet illos de terra sua 6:2 locutusque est Do-

the LORD: 6:3 And I appeared unto Abraham, unto Isaac, and unto minus ad Mosen dicens ego Dominus 6:3 qui apparui Ab-Jacob, by the name of God Almighty, but by my name JEHOVAH was I not known to them.

raham Isaac et Iacob in Deo omnipotente et nomen meum

6:4 And I have also established my covenant with them, to Adonai non indicaui eis 6:4 pepigique cum eis foedus ut give them the land of Canaan, the land of their pilgrimage, wherein they were strangers.

darem illis terram Chanaan terram peregrinationis eorum in

6:5 And I have also heard the groaning of the children of qua fuerunt aduenae 6:5 ego audiui gemitum filiorum Isra-

Israel, whom the Egyptians keep in bondage; and I have remembered my covenant.

hel quo Aegyptii oppresserunt eos et recordatus sum pacti

6:6 Wherefore say unto the children of Israel, I am the mei 6:6 ideo dic filiis Israhel ego Dominus qui educam uos

LORD, and I will bring you out from under the burdens of the Egyptians, and I will rid you out of their bondage, and I will redeem you with a stretched out arm, and with great judgments:

de ergastulo Aegyptiorum et eruam de seruitute ac rediiam

6:7 And I will take you to me for a people, and I will be in brachio excelso et iudiciis magnis 6:7 et adsumam uos

to you a God: and ye shall know that I am the LORD your God, which bringeth you out from under the burdens of

the Egyptians. mihi in populum et ero uester Deus scietisque quod ego sim

Dominus Deus uester qui eduxerim uos de ergastulo Ae-

6:8 And I will bring you in unto the land, concerning the gyptiorum 6:8 et induxerim in terram super quam leuaui which I did swear to give it to Abraham, to Isaac, and to Jacob; and I will give it you for an heritage: I am the

LORD. manum meam ut darem eam Abraham Isaac et Iacob da-6:9 And Moses spake so unto the children of Israel: but boque illam uobis possidendam ego Dominus 6:9 narrauit

they hearkened not unto Moses for anguish of spirit, and for cruel bondage.

ergo Moses omnia filiis Israhel qui non adquieuerunt ei pro-

6:10 And the LORD spake unto Moses, saying, pter angustiam spiritus et opus durissimum 6:10 locutusque 6:11 Go in, speak unto Pharaoh king of Egypt, that he let est Dominus ad Mosen dicens 6:11 ingredere et loquere ad

the children of Israel go out of his land.

Pharao regem Aegypti ut dimittat filios Israhel de terra sua 6:12 respondit Moses coram Domino ecce filii Israhel non 6:12 And Moses spake before the LORD, saying, Behold,

the children of Israel have not hearkened unto me; how then shall Pharaoh hear me, who am of uncircumcised

me audiunt et quomodo audiet me Pharao praesertim cum lips? sim incircumcisus labiis 6:13 locutus est Dominus ad Mo- 6:13 And the LORD spake unto Moses and unto Aaron,

and gave them a charge unto the children of Israel, and unto Pharaoh king of Egypt, to bring the children of Israel

sen et Aaron et dedit mandatum ad filios Israhel et ad Pharao out of the land of Egypt. regem Aegypti ut educerent filios Israhel de terra Aegypti 6:14 isti sunt principes domorum per familias suas filii Ru- 6:14 These be the heads of their fathers’ houses: The

sons of Reuben the firstborn of Israel; Hanoch, and Pallu, Hezron, and Carmi: these be the families of Reuben.

ben primogeniti Israhelis Enoch et Phallu Aesrom et Charmi 6:15 hae cognationes Ruben filii Symeon Iamuhel et Iamin et 6:15 And the sons of Simeon; Jemuel, and Jamin, and

Ohad, and Jachin, and Zohar, and Shaul the son of a Canaanitish woman: these are the families of Simeon.

Aod Iachin et Soer et Saul filius Chananitidis hae progenies Symeon 6:16 et haec nomina filiorum Leui per cognationes 6:16 And these are the names of the sons of Levi according

to their generations; Gershon, and Kohath, and Merari: and the years of the life of Levi were an hundred thirty

suas Gerson et Caath et Merari anni autem uitae Leui fuerunt and seven years. centum triginta septem 6:17 filii Gerson Lobeni et Semei 6:17 The sons of Gershon; Libni, and Shimi, according to

their families.

per cognationes suas 6:18 filii Caath Amram et Isuar et He- 6:18 And the sons of Kohath; Amram, and Izhar, and He-

bron, and Uzziel: and the years of the life of Kohath were an hundred thirty and three years.

bron et Ozihel annique uitae Caath centum triginta tres 6:19

6:19 And the sons of Merari; Mahali and Mushi: these are the families of Levi according to their generations.

filii Merari Mooli et Musi hae cognationes Leui per familias suas 6:20 accepit autem Amram uxorem Iocabed patruelem 6:20 And Amram took him Jochebed his father’s sister to

wife; and she bare him Aaron and Moses: and the years of the life of Amram were an hundred and thirty and seven

suam quae peperit ei Aaron et Mosen fueruntque anni ui- years. tae Amram centum triginta septem 6:21 filii quoque Isuar 6:21 And the sons of Izhar; Korah, and Nepheg, and

Zichri.

Core et Napheg et Zechri 6:22 filii quoque Ozihel Misa- 6:22 And the sons of Uzziel; Mishael, and Elzaphan, and

Zithri.

hel et Elsaphan et Sethri 6:23 accepit autem Aaron uxorem 6:23 And Aaron took him Elisheba, daughter of Ammi-

nadab, sister of Naashon, to wife; and she bare him Nadab, and Abihu, Eleazar, and Ithamar.

Elisabe filiam Aminadab sororem Naasson quae peperit ei Nadab et Abiu et Eleazar et Ithamar 6:24 filii quoque Core 6:24 And the sons of Korah; Assir, and Elkanah, and Abi-

asaph: these are the families of the Korhites.

Asir et Helcana et Abiasab hae sunt cognationes Coritarum 6:25 at uero Eleazar filius Aaron accepit uxorem de filiabus 6:25 And Eleazar Aaron’s son took him one of the daugh-

ters of Putiel to wife; and she bare him Phinehas: these are the heads of the fathers of the Levites according to

Phutihel quae peperit ei Finees hii sunt principes familiarum their families. leuiticarum per cognationes suas 6:26 iste est Aaron et Mo- 6:26 These are that Aaron and Moses, to whom the LORD

said, Bring out the children of Israel from the land of Egypt according to their armies.

ses quibus praecepit Dominus ut educerent filios Israhel de terra Aegypti per turmas suas 6:27 hii sunt qui loquuntur ad 6:27 These are they which spake to Pharaoh king of Egypt,

to bring out the children of Israel from Egypt: these are that Moses and Aaron.

Pharao regem Aegypti ut educant filios Israhel de Aegypto iste Moses et Aaron 6:28 in die qua locutus est Dominus ad 6:28 And it came to pass on the day when the LORD spake

unto Moses in the land of Egypt,

Mosen in terra Aegypti 6:29 et locutus est Dominus ad Mo- 6:29 That the LORD spake unto Moses, saying, I am the

LORD: speak thou unto Pharaoh king of Egypt all that I say unto thee.

sen dicens ego Dominus loquere ad Pharao regem Aegypti omnia quae ego loquor tibi 6:30 et ait Moses coram Do- 6:30 And Moses said before the LORD, Behold, I am of

uncircumcised lips, and how shall Pharaoh hearken unto me?

Biblia Sacra Vulgata

mino en incircumcisus labiis sum quomodo audiet me Pharao

7:1 And the LORD said unto Moses, See, I have made 7:1 dixitque Dominus ad Mosen ecce constitui te Deum

thee a god to Pharaoh: and Aaron thy brother shall be thy prophet.

Pharaonis Aaron frater tuus erit propheta tuus 7:2 tu loque-

7:2 Thou shalt speak all that I command thee: and Aaron thy brother shall speak unto Pharaoh, that he send the children of Israel out of his land.

ris omnia quae mando tibi ille loquetur ad Pharaonem ut di-

7:3 And I will harden Pharaoh’s heart, and multiply my mittat filios Israhel de terra sua 7:3 sed ego indurabo cor

signs and my wonders in the land of Egypt.

eius et multiplicabo signa et ostenta mea in terra Aegypti

7:4 But Pharaoh shall not hearken unto you, that I may lay 7:4 et non audiet uos inmittamque manum meam super Aemy hand upon Egypt, and bring forth mine armies, and my people the children of Israel, out of the land of Egypt by

great judgments. gyptum et educam exercitum et populum meum filios Isra-7:5 And the Egyptians shall know that I am the LORD, hel de terra Aegypti per iudicia maxima 7:5 et scient Ae-

when I stretch forth mine hand upon Egypt, and bring out the children of Israel from among them.

gyptii quod ego sim Dominus qui extenderim manum meam super Aegyptum et eduxerim filios Israhel de medio eorum

7:6 And Moses and Aaron did as the LORD commanded 7:6 fecit itaque Moses et Aaron sicut praeceperat Dominus

them, so did they.

7:7 And Moses was fourscore years old, and Aaron ita egerunt 7:7 erat autem Moses octoginta annorum et Aa-

fourscore and three years old, when they spake unto Pharaoh.

ron octoginta trium quando locuti sunt ad Pharaonem 7:8

7:8 And the LORD spake unto Moses and unto Aaron, saying,

7:9 When Pharaoh shall speak unto you, saying, Shew a dixitque Dominus ad Mosen et Aaron 7:9 cum dixerit uo-

miracle for you: then thou shalt say unto Aaron, Take thy rod, and cast it before Pharaoh, and it shall become a ser-

pent. bis Pharao ostendite signa dices ad Aaron tolle uirgam tuam 7:10 And Moses and Aaron went in unto Pharaoh, and et proice eam coram Pharao ac uertatur in colubrum 7:10

they did so as the LORD had commanded: and Aaron cast down his rod before Pharaoh, and before his servants, and

it became a serpent. ingressi itaque Moses et Aaron ad Pharaonem fecerunt sicut

praeceperat Dominus tulitque Aaron uirgam coram Pharao

7:11 Then Pharaoh also called the wise men and the sor- et seruis eius quae uersa est in colubrum 7:11 uocauit autem

cerers: now the magicians of Egypt, they also did in like manner with their enchantments.

Pharao sapientes et maleficos et fecerunt etiam ipsi per in-

7:12 For they cast down every man his rod, and they be- cantationes aegyptias et arcana quaedam similiter 7:12 pro-

came serpents: but Aaron’s rod swallowed up their rods.

ieceruntque singuli uirgas suas quae uersae sunt in dracones

7:13 And he hardened Pharaoh’s heart, that he hearkened sed deuorauit uirga Aaron uirgas eorum 7:13 induratumque

not unto them; as the LORD had said.

est cor Pharaonis et non audiuit eos sicut praeceperat Domi-

7:14 And the LORD said unto Moses, Pharaoh’s heart is nus 7:14 dixit autem Dominus ad Mosen ingrauatum est cor

hardened, he refuseth to let the people go. 7:15 Get thee unto Pharaoh in the morning; lo, he goeth Pharaonis non uult dimittere populum 7:15 uade ad eum out unto the water; and thou shalt stand by the river’s brink against he come; and the rod which was turned to a serpent

shalt thou take in thine hand. mane ecce egredietur ad aquas et stabis in occursum eius su-

per ripam fluminis et uirgam quae conuersa est in draconem

7:16 And thou shalt say unto him, The LORD God of the tolles in manu tua 7:16 dicesque ad eum Dominus Deus He-

Hebrews hath sent me unto thee, saying, Let my people go, that they may serve me in the wilderness: and, behold,

hitherto thou wouldest not hear. braeorum misit me ad te dicens dimitte populum meum ut mihi sacrificet in deserto et usque ad praesens audire noluisti 7:17 haec igitur dicit Dominus in hoc scies quod Dominus 7:17 Thus saith the LORD, In this thou shalt know that I

am the LORD: behold, I will smite with the rod that is in mine hand upon the waters which are in the river, and they

sim ecce percutiam uirga quae in manu mea est aquam flu- shall be turned to blood. minis et uertetur in sanguinem 7:18 pisces quoque qui sunt 7:18 And the fish that is in the river shall die, and the river

shall stink; and the Egyptians shall lothe to drink of the water of the river.

in fluuio morientur et conputrescent aquae et adfligentur Aegyptii bibentes aquam fluminis 7:19 dixit quoque Dominus 7:19 And the LORD spake unto Moses, Say unto Aaron,

Take thy rod, and stretch out thine hand upon the waters of Egypt, upon their streams, upon their rivers, and upon their ponds, and upon all their pools of water, that they may become blood; and that there may be blood throughout all the land of Egypt, both in vessels of wood, and in vessels of stone.

ad Mosen dic ad Aaron tolle uirgam tuam et extende manum tuam super aquas Aegypti et super fluuios eorum et riuos ac paludes et omnes lacus aquarum ut uertantur in sanguinem et sit cruor in omni terra Aegypti tam in ligneis uasis quam in saxeis 7:20 feceruntque ita Moses et Aaron sicut prae- 7:20 And Moses and Aaron did so, as the LORD com-

manded; and he lifted up the rod, and smote the waters that were in the river, in the sight of Pharaoh, and in the sight of his servants; and all the waters that were in the river were turned to blood.

ceperat Dominus et eleuans uirgam percussit aquam fluminis coram Pharao et seruis eius quae uersa est in sanguinem 7:21 et pisces qui erant in flumine mortui sunt conputruitque 7:21 And the fish that was in the river died; and the river

stank, and the Egyptians could not drink of the water of the river; and there was blood throughout all the land of

fluuius et non poterant Aegyptii bibere aquam fluminis et Egypt. fuit sanguis in tota terra Aegypti 7:22 feceruntque simili- 7:22 And the magicians of Egypt did so with their enchant-

ments: and Pharaoh’s heart was hardened, neither did he hearken unto them; as the LORD had said.

ter malefici Aegyptiorum incantationibus suis et induratum est cor Pharaonis nec audiuit eos sicut praeceperat Dominus 7:23 auertitque se et ingressus est domum suam nec adpo- 7:23 And Pharaoh turned and went into his house, neither

did he set his heart to this also.

suit cor etiam hac uice 7:24 foderunt autem omnes Aegyptii 7:24 And all the Egyptians digged round about the river

for water to drink; for they could not drink of the water of the river.

per circuitum fluminis aquam ut biberent non enim poterant bibere de aqua fluminis 7:25 impletique sunt septem dies 7:25 And seven days were fulfilled, after that the LORD

had smitten the river.

postquam percussit Dominus fluuium

8:1 dixitque Dominus ad Mosen ingredere ad Pharao et 8:1 And the LORD spake unto Moses, Go unto Pharaoh,

and say unto him, Thus saith the LORD, Let my people go, that they may serve me.

dices ad eum haec dicit Dominus dimitte populum meum ut sacrificet mihi 8:2 sin autem nolueris dimittere ecce ego 8:2 And if thou refuse to let them go, behold, I will smite

all thy borders with frogs:

percutiam omnes terminos tuos ranis 8:3 et ebulliet fluuius 8:3 And the river shall bring forth frogs abundantly, which

shall go up and come into thine house, and into thy bedchamber, and upon thy bed, and into the house of thy servants, and upon thy people, and into thine ovens, and into thy kneadingtroughs:

ranas quae ascendent et ingredientur domum tuam et cubiculum lectuli tui et super stratum tuum et in domos seruorum tuorum et in populum tuum et in furnos tuos et in reliquias ciborum tuorum 8:4 et ad te et ad populum tuum et ad om- 8:4 And the frogs shall come up both on thee, and upon

thy people, and upon all thy servants.

nes seruos tuos intrabunt ranae 8:5 dixitque Dominus ad 8:5 And the LORD spake unto Moses, Say unto Aaron,

Stretch forth thine hand with thy rod over the streams, over the rivers, and over the ponds, and cause frogs to come up

Mosen dic Aaron extende manum tuam super fluuios et su- upon the land of Egypt. per riuos ac paludes et educ ranas super terram Aegypti 8:6 8:6 And Aaron stretched out his hand over the waters of

Egypt; and the frogs came up, and covered the land of Egypt.

Biblia Sacra Vulgata

extendit Aaron manum super aquas Aegypti et ascenderunt

8:7 And the magicians did so with their enchantments, and ranae operueruntque terram Aegypti 8:7 fecerunt autem et

brought up frogs upon the land of Egypt.

malefici per incantationes suas similiter eduxeruntque ranas

8:8 Then Pharaoh called for Moses and Aaron, and said, super terram Aegypti 8:8 uocauit autem Pharao Mosen et

Intreat the LORD, that he may take away the frogs from me, and from my people; and I will let the people go, that

they may do sacrifice unto the LORD. Aaron et dixit orate Dominum ut auferat ranas a me et a 8:9 And Moses said unto Pharaoh, Glory over me: when populo meo et dimittam populum ut sacrificet Domino 8:9

shall I intreat for thee, and for thy servants, and for thy people, to destroy the frogs from thee and thy houses, that

they may remain in the river only? dixitque Moses Pharaoni constitue mihi quando deprecer pro

te et pro seruis tuis et pro populo tuo ut abigantur ranae a

8:10 And he said, To morrow. And he said, Be it according te et a domo tua et tantum in flumine remaneant 8:10 qui to thy word: that thou mayest know that there is none like unto the LORD our God.

respondit cras at ille iuxta uerbum inquit tuum ut scias quo-

8:11 And the frogs shall depart from thee, and from thy niam non est sicut Dominus Deus noster 8:11 et recedent

houses, and from thy servants, and from thy people; they shall remain in the river only.

ranae a te et a domo tua et a seruis tuis et a populo tuo tan-

8:12 And Moses and Aaron went out from Pharaoh: and tum in flumine remanebunt 8:12 egressique sunt Moses et

Moses cried unto the LORD because of the frogs which he had brought against Pharaoh.

Aaron a Pharaone et clamauit Moses ad Dominum pro spon-

8:13 And the LORD did according to the word of Moses; sione ranarum quam condixerat Pharaoni 8:13 fecitque Do-

and the frogs died out of the houses, out of the villages, and out of the fields.

minus iuxta uerbum Mosi et mortuae sunt ranae de domibus

8:14 And they gathered them together upon heaps: and the et de uillis et de agris 8:14 congregaueruntque eas in in-

land stank.

8:15 But when Pharaoh saw that there was respite, he mensos aggeres et conputruit terra 8:15 uidens autem Pha-

hardened his heart, and hearkened not unto them; as the LORD had said.

rao quod data esset requies ingrauauit cor suum et non audi-

8:16 And the LORD said unto Moses, Say unto Aaron, uit eos sicut praeceperat Dominus 8:16 dixitque Dominus

Stretch out thy rod, and smite the dust of the land, that it may become lice throughout all the land of Egypt.

ad Mosen loquere ad Aaron extende uirgam tuam et percute puluerem terrae et sint scinifes in uniuersa terra Aegypti

8:17 And they did so; for Aaron stretched out his hand 8:17 feceruntque ita et extendit Aaron manu uirgam tenens

with his rod, and smote the dust of the earth, and it became lice in man, and in beast; all the dust of the land became

lice throughout all the land of Egypt. percussitque puluerem terrae et facti sunt scinifes in homi-

nibus et in iumentis omnis puluis terrae uersus est in sci-

8:18 And the magicians did so with their enchantments to nifes per totam terram Aegypti 8:18 feceruntque similiter

bring forth lice, but they could not: so there were lice upon man, and upon beast.

malefici incantationibus suis ut educerent scinifes et non potuerunt erantque scinifes tam in hominibus quam in iumentis

8:19 Then the magicians said unto Pharaoh, This is the 8:19 et dixerunt malefici ad Pharao digitus Dei est induratum-

finger of God: and Pharaoh’s heart was hardened, and he hearkened not unto them; as the LORD had said.

que est cor Pharaonis et non audiuit eos sicut praeceperat

8:20 And the LORD said unto Moses, Rise up early in the Dominus 8:20 dixit quoque Dominus ad Mosen consurge morning, and stand before Pharaoh; lo, he cometh forth to the water; and say unto him, Thus saith the LORD, Let my

people go, that they may serve me. diluculo et sta coram Pharaone egreditur enim ad aquas et dices ad eum haec dicit Dominus dimitte populum meum ut sacrificet mihi 8:21 quod si non dimiseris eum ecce ego in- 8:21 Else, if thou wilt not let my people go, behold, I will

send swarms of flies upon thee, and upon thy servants, and upon thy people, and into thy houses: and the houses of the Egyptians shall be full of swarms of flies, and also the ground whereon they are.

mittam in te et in seruos tuos et in populum tuum et in domos tuas omne genus muscarum et implebuntur domus Aegyptiorum muscis diuersi generis et in uniuersa terra in qua fuerint 8:22 faciamque mirabilem in die illa terram Gessen in 8:22 And I will sever in that day the land of Goshen, in

which my people dwell, that no swarms of flies shall be there; to the end thou mayest know that I am the LORD in

qua populus meus est ut non sint ibi muscae et scias quoniam the midst of the earth. ego Dominus in medio terrae 8:23 ponamque diuisionem 8:23 And I will put a division between my people and thy

people: to morrow shall this sign be.

inter populum meum et populum tuum cras erit signum istud 8:24 fecitque Dominus ita et uenit musca grauissima in do- 8:24 And the LORD did so; and there came a grievous

swarm of flies into the house of Pharaoh, and into his servants’ houses, and into all the land of Egypt: the land was

mos Pharaonis et seruorum eius et in omnem terram Aegypti corrupted by reason of the swarm of flies. corruptaque est terra ab huiuscemodi muscis 8:25 uocauit 8:25 And Pharaoh called for Moses and for Aaron, and

said, Go ye, sacrifice to your God in the land.

Pharao Mosen et Aaron et ait eis ite sacrificate Deo uestro in terra 8:26 et ait Moses non potest ita fieri abominationes 8:26 And Moses said, It is not meet so to do; for we shall

sacrifice the abomination of the Egyptians to the LORD our God: lo, shall we sacrifice the abomination of the

enim Aegyptiorum immolabimus Domino Deo nostro quod Egyptians before their eyes, and will they not stone us?

si mactauerimus ea quae colunt Aegyptii coram eis lapidibus nos obruent 8:27 uia trium dierum pergemus in solitudine 8:27 We will go three days’ journey into the wilderness,

and sacrifice to the LORD our God, as he shall command us.

et sacrificabimus Domino Deo nostro sicut praeceperit nobis 8:28 dixitque Pharao ego dimittam uos ut sacrificetis Domino 8:28 And Pharaoh said, I will let you go, that ye may sac-

rifice to the LORD your God in the wilderness; only ye shall not go very far away: intreat for me.

Deo uestro in deserto uerumtamen longius ne abeatis rogate pro me 8:29 et ait Moses egressus a te orabo Dominum et 8:29 And Moses said, Behold, I go out from thee, and I

will intreat the LORD that the swarms of flies may depart from Pharaoh, from his servants, and from his people, to morrow: but let not Pharaoh deal deceitfully any more in not letting the people go to sacrifice to the LORD.

recedet musca a Pharaone et a seruis et a populo eius cras uerumtamen noli ultra fallere ut non dimittas populum sacrificare Domino 8:30 egressusque Moses a Pharao orauit 8:30 And Moses went out from Pharaoh, and intreated the

LORD.

Dominum 8:31 qui fecit iuxta uerbum illius et abstulit mu- 8:31 And the LORD did according to the word of Moses;

and he removed the swarms of flies from Pharaoh, from his servants, and from his people; there remained not one.

scas a Pharao et a seruis et a populo eius non superfuit ne

una quidem 8:32 et ingrauatum est cor Pharaonis ita ut ne 8:32 And Pharaoh hardened his heart at this time also, nei-

ther would he let the people go.

hac quidem uice dimitteret populum

9:1 dixit autem Dominus ad Mosen ingredere ad Pharao- 9:1 Then the LORD said unto Moses, Go in unto Pharaoh,

and tell him, Thus saith the LORD God of the Hebrews, Let my people go, that they may serve me.

nem et loquere ad eum haec dicit Dominus Deus Hebraeorum dimitte populum meum ut sacrificet mihi 9:2 quod si 9:2 For if thou refuse to let them go, and wilt hold them

still,

adhuc rennuis et retines eos 9:3 ecce manus mea erit su- 9:3 Behold, the hand of the LORD is upon thy cattle which

is in the field, upon the horses, upon the asses, upon the camels, upon the oxen, and upon the sheep: there shall be

per agros tuos et super equos et asinos et camelos et boues a very grievous murrain. et oues pestis ualde grauis 9:4 et faciet Dominus mirabile 9:4 And the LORD shall sever between the cattle of Israel

and the cattle of Egypt: and there shall nothing die of all that is the children’s of Israel.

Biblia Sacra Vulgata

inter possessiones Israhel et possessiones Aegyptiorum ut nihil omnino intereat ex his quae pertinent ad filios Israhel

9:5 And the LORD appointed a set time, saying, To mor- 9:5 constituitque Dominus tempus dicens cras faciet Domi-

row the LORD shall do this thing in the land.

9:6 And the LORD did that thing on the morrow, and all nus uerbum istud in terra 9:6 fecit ergo Dominus uerbum

the cattle of Egypt died: but of the cattle of the children of Israel died not one.

hoc altero die mortuaque sunt omnia animantia Aegyptiorum de animalibus uero filiorum Israhel nihil omnino periit

9:7 And Pharaoh sent, and, behold, there was not one of 9:7 et misit Pharao ad uidendum nec erat quicquam mortuum

the cattle of the Israelites dead. And the heart of Pharaoh was hardened, and he did not let the people go.

de his quae possidebat Israhel ingrauatumque est cor Phara-

9:8 And the LORD said unto Moses and unto Aaron, Take onis et non dimisit populum 9:8 et dixit Dominus ad Mosen

to you handfuls of ashes of the furnace, and let Moses sprinkle it toward the heaven in the sight of Pharaoh.

et Aaron tollite plenas manus cineris de camino et spargat il-

9:9 And it shall become small dust in all the land of Egypt, lud Moses in caelum coram Pharao 9:9 sitque puluis super and shall be a boil breaking forth with blains upon man, and upon beast, throughout all the land of Egypt.

omnem terram Aegypti erunt enim in hominibus et in iumentis uulnera et uesicae turgentes in uniuersa terra Aegypti

9:10 And they took ashes of the furnace, and stood before 9:10 tuleruntque cinerem de camino et steterunt contra Pha-

Pharaoh; and Moses sprinkled it up toward heaven; and it became a boil breaking forth with blains upon man, and

upon beast. rao et sparsit illud Moses in caelum factaque sunt uulnera 9:11 And the magicians could not stand before Moses be- uesicarum turgentium in hominibus et in iumentis 9:11 nec

cause of the boils; for the boil was upon the magicians, and upon all the Egyptians.

poterant malefici stare coram Mosen propter uulnera quae in

9:12 And the LORD hardened the heart of Pharaoh, and he illis erant et in omni terra Aegypti 9:12 indurauitque Domi-

hearkened not unto them; as the LORD had spoken unto Moses.

nus cor Pharaonis et non audiuit eos sicut locutus est Domi-

9:13 And the LORD said unto Moses, Rise up early in nus ad Mosen 9:13 dixit quoque Dominus ad Mosen mane

the morning, and stand before Pharaoh, and say unto him, Thus saith the LORD God of the Hebrews, Let my people

go, that they may serve me. consurge et sta coram Pharao et dices ad eum haec dicit Do-

minus Deus Hebraeorum dimitte populum meum ut sacrifi-

9:14 For I will at this time send all my plagues upon thine cet mihi 9:14 quia in hac uice mittam omnes plagas meas heart, and upon thy servants, and upon thy people; that thou mayest know that there is none like me in all the

earth. super cor tuum super seruos tuos et super populum tuum ut 9:15 For now I will stretch out my hand, that I may smite scias quod non sit similis mei in omni terra 9:15 nunc enim

thee and thy people with pestilence; and thou shalt be cut off from the earth.

extendens manum percutiam te et populum tuum peste per-

9:16 And in very deed for this cause have I raised thee up, ibisque de terra 9:16 idcirco autem posui te ut ostendam in

for to shew in thee my power; and that my name may be declared throughout all the earth.

te fortitudinem meam et narretur nomen meum in omni terra

9:17 As yet exaltest thou thyself against my people, that 9:17 adhuc retines populum meum et non uis eum dimittere thou wilt not let them go? 9:18 Behold, to morrow about this time I will cause it to 9:18 en pluam hac ipsa hora cras grandinem multam nimis rain a very grievous hail, such as hath not been in Egypt since the foundation thereof even until now.

qualis non fuit in Aegypto a die qua fundata est usque in

9:19 Send therefore now, and gather thy cattle, and all that praesens tempus 9:19 mitte ergo iam nunc et congrega iuthou hast in the field; for upon every man and beast which shall be found in the field, and shall not be brought home,

the hail shall come down upon them, and they shall die. menta tua et omnia quae habes in agro homines enim et iumenta et uniuersa quae inuenta fuerint foris nec congregata de agris cecideritque super ea grando morientur 9:20 qui ti- 9:20 He that feared the word of the LORD among the ser-

vants of Pharaoh made his servants and his cattle flee into the houses:

muit uerbum Domini de seruis Pharao fecit confugere seruos suos et iumenta in domos 9:21 qui autem neglexit sermo- 9:21 And he that regarded not the word of the LORD left

his servants and his cattle in the field.

nem Domini dimisit seruos suos et iumenta in agris 9:22 9:22 And the LORD said unto Moses, Stretch forth thine

hand toward heaven, that there may be hail in all the land of Egypt, upon man, and upon beast, and upon every herb

et dixit Dominus ad Mosen extende manum tuam in cae- of the field, throughout the land of Egypt. lum ut fiat grando in uniuersa terra Aegypti super homines et super iumenta et super omnem herbam agri in terra Aegypti 9:23 extenditque Moses uirgam in caelum et Domi- 9:23 And Moses stretched forth his rod toward heaven:

and the LORD sent thunder and hail, and the fire ran along upon the ground; and the LORD rained hail upon the land

nus dedit tonitrua et grandinem ac discurrentia fulgura super of Egypt. terram pluitque Dominus grandinem super terram Aegypti 9:24 et grando et ignis inmixta pariter ferebantur tantaeque 9:24 So there was hail, and fire mingled with the hail, very

grievous, such as there was none like it in all the land of Egypt since it became a nation.

fuit magnitudinis quanta ante numquam apparuit in uniuersa terra Aegypti ex quo gens illa condita est 9:25 et percussit 9:25 And the hail smote throughout all the land of Egypt

all that was in the field, both man and beast; and the hail smote every herb of the field, and brake every tree of the

grando in omni terra Aegypti cuncta quae fuerunt in agris field. ab homine usque ad iumentum cunctam herbam agri percussit grando et omne lignum regionis confregit 9:26 tan- 9:26 Only in the land of Goshen, where the children of

Israel were, was there no hail.

tum in terra Gessen ubi erant filii Israhel grando non cecidit 9:27 misitque Pharao et uocauit Mosen et Aaron dicens 9:27 And Pharaoh sent, and called for Moses and Aaron,

and said unto them, I have sinned this time: the LORD is righteous, and I and my people are wicked.

ad eos peccaui etiam nunc Dominus iustus ego et populus meus impii 9:28 orate Dominum et desinant tonitrua Dei 9:28 Intreat the LORD (for it is enough) that there be no

more mighty thunderings and hail; and I will let you go, and ye shall stay no longer.

et grando ut dimittam uos et nequaquam hic ultra maneatis 9:29 ait Moses cum egressus fuero de urbe extendam palmas 9:29 And Moses said unto him, As soon as I am gone

out of the city, I will spread abroad my hands unto the LORD; and the thunder shall cease, neither shall there be any more hail; that thou mayest know how that the earth is the LORD’s.

meas ad Dominum et cessabunt tonitrua et grando non erit ut scias quia Domini est terra 9:30 noui autem quod et tu 9:30 But as for thee and thy servants, I know that ye will

not yet fear the LORD God.

et serui tui necdum timeatis Dominum Deum 9:31 linum 9:31 And the flax and the barley was smitten: for the bar-

ley was in the ear, and the flax was bolled.

ergo et hordeum laesum est eo quod hordeum esset uirens et linum iam folliculos germinaret 9:32 triticum autem et far 9:32 But the wheat and the rie were not smitten: for they

were not grown up.

non sunt laesa quia serotina erant 9:33 egressusque Moses a 9:33 And Moses went out of the city from Pharaoh, and

spread abroad his hands unto the LORD: and the thunders and hail ceased, and the rain was not poured upon the

Pharaone et ex urbe tetendit manus ad Dominum et cessaue- earth. runt tonitrua et grando nec ultra stillauit pluuia super terram 9:34 uidens autem Pharao quod cessasset pluuia et grando et 9:34 And when Pharaoh saw that the rain and the hail and

the thunders were ceased, he sinned yet more, and hardened his heart, he and his servants.

tonitrua auxit peccatum 9:35 et ingrauatum est cor eius et

9:35 And the heart of Pharaoh was hardened, neither would he let the children of Israel go; as the LORD had spoken by Moses.

Biblia Sacra Vulgata

seruorum illius et induratum nimis nec dimisit filios Israhel sicut praeceperat Dominus per manum Mosi

10:1 And the LORD said unto Moses, Go in unto Pharaoh: 10:1 et dixit Dominus ad Mosen ingredere ad Pharao ego

for I have hardened his heart, and the heart of his servants, that I might shew these my signs before him:

enim induraui cor eius et seruorum illius ut faciam signa mea

10:2 And that thou mayest tell in the ears of thy son, and haec in eo 10:2 et narres in auribus filii tui et nepotum tuo-

of thy son’s son, what things I have wrought in Egypt, and my signs which I have done among them; that ye may

know how that I am the LORD. rum quotiens contriuerim Aegyptios et signa mea fecerim in 10:3 And Moses and Aaron came in unto Pharaoh, and eis et sciatis quia ego Dominus 10:3 introierunt ergo Mo-

said unto him, Thus saith the LORD God of the Hebrews, How long wilt thou refuse to humble thyself before me?

let my people go, that they may serve me. ses et Aaron ad Pharaonem et dixerunt ad eum haec dicit

Dominus Deus Hebraeorum usquequo non uis subici mihi

10:4 Else, if thou refuse to let my people go, behold, to dimitte populum meum ut sacrificet mihi 10:4 sin autem morrow will I bring the locusts into thy coast:

resistis et non uis dimittere eum ecce ego inducam cras lu-

10:5 And they shall cover the face of the earth, that one custam in fines tuos 10:5 quae operiat superficiem terrae

cannot be able to see the earth: and they shall eat the residue of that which is escaped, which remaineth unto you from the hail, and shall eat every tree which groweth for you out of the field:

nec quicquam eius appareat sed comedatur quod residuum fuit grandini conrodet enim omnia ligna quae germinant in

10:6 And they shall fill thy houses, and the houses of all agris 10:6 et implebunt domos tuas et seruorum tuorum et

thy servants, and the houses of all the Egyptians; which neither thy fathers, nor thy fathers’ fathers have seen, since the day that they were upon the earth unto this day. And he turned himself, and went out from Pharaoh.

omnium Aegyptiorum quantam non uiderunt patres tui et aui ex quo orti sunt super terram usque in praesentem diem

10:7 And Pharaoh’s servants said unto him, How long auertitque se et egressus est a Pharaone 10:7 dixerunt au-

shall this man be a snare unto us? let the men go, that they may serve the LORD their God: knowest thou not

yet that Egypt is destroyed? tem serui Pharaonis ad eum usquequo patiemur hoc scanda-

lum dimitte homines ut sacrificent Domino Deo suo nonne

10:8 And Moses and Aaron were brought again unto uides quod perierit Aegyptus 10:8 reuocaueruntque Mosen

Pharaoh: and he said unto them, Go, serve the LORD your God: but who are they that shall go?

et Aaron ad Pharaonem qui dixit eis ite sacrificate Domino

10:9 And Moses said, We will go with our young and with Deo uestro quinam sunt qui ituri sunt 10:9 ait Moses cum

our old, with our sons and with our daughters, with our flocks and with our herds will we go; for we must hold a

feast unto the LORD. paruulis nostris et senibus pergemus cum filiis et filiabus

cum ouibus et armentis est enim sollemnitas Domini nos-

10:10 And he said unto them, Let the LORD be so with tri 10:10 et respondit sic Dominus sit uobiscum quomodo

you, as I will let you go, and your little ones: look to it; for evil is before you.

ego dimittam uos et paruulos uestros cui dubium est quod

10:11 Not so: go now ye that are men, and serve the pessime cogitetis 10:11 non fiet ita sed ite tantum uiri et sa-

LORD; for that ye did desire. And they were driven out from Pharaoh’s presence.

crificate Domino hoc enim et ipsi petistis statimque eiecti

10:12 And the LORD said unto Moses, Stretch out thine sunt de conspectu Pharaonis 10:12 dixit autem Dominus ad

hand over the land of Egypt for the locusts, that they may come up upon the land of Egypt, and eat every herb of the

land, even all that the hail hath left. Mosen extende manum tuam super terram Aegypti ad lu-

custam ut ascendat super eam et deuoret omnem herbam

10:13 And Moses stretched forth his rod over the land of quae residua fuit grandini 10:13 extendit Moses uirgam su-

Egypt, and the LORD brought an east wind upon the land all that day, and all that night; and when it was morning, the east wind brought the locusts.

per terram Aegypti et Dominus induxit uentum urentem tota illa die ac nocte et mane facto uentus urens leuauit lucustas 10:14 quae ascenderunt super uniuersam terram Aegypti et 10:14 And the locust went up over all the land of Egypt,

and rested in all the coasts of Egypt: very grievous were they; before them there were no such locusts as they, nei-

sederunt in cunctis finibus Aegyptiorum innumerabiles qua- ther after them shall be such. les ante illud tempus non fuerant nec postea futurae sunt 10:15 operueruntque uniuersam superficiem terrae uastantes 10:15 For they covered the face of the whole earth, so that

the land was darkened; and they did eat every herb of the land, and all the fruit of the trees which the hail had left: and there remained not any green thing in the trees, or in the herbs of the field, through all the land of Egypt.

omnia deuorata est igitur herba terrae et quicquid pomorum in arboribus fuit quae grando dimiserat nihilque omnino uirens relictum est in lignis et in herbis terrae in cuncta Ae-

gypto 10:16 quam ob rem festinus Pharao uocauit Mosen 10:16 Then Pharaoh called for Moses and Aaron in haste; and he said, I have sinned against the LORD your God, and against you.

et Aaron et dixit eis peccaui in Dominum Deum uestrum et in uos 10:17 sed nunc dimittite peccatum mihi etiam hac 10:17 Now therefore forgive, I pray thee, my sin only this

once, and intreat the LORD your God, that he may take away from me this death only.

uice et rogate Dominum Deum uestrum ut auferat a me mortem istam 10:18 egressusque est de conspectu Pharaonis et 10:18 And he went out from Pharaoh, and intreated the

LORD.

orauit Dominum 10:19 qui flare fecit uentum ab occidente 10:19 And the LORD turned a mighty strong west wind,

which took away the locusts, and cast them into the Red sea; there remained not one locust in all the coasts of

uehementissimum et arreptam lucustam proiecit in mare Ru- Egypt. brum non remansit ne una quidem in cunctis finibus Aegypti 10:20 et indurauit Dominus cor Pharaonis nec dimisit filios 10:20 But the LORD hardened Pharaoh’s heart, so that he

would not let the children of Israel go.

Israhel 10:21 dixit autem Dominus ad Mosen extende ma- 10:21 And the LORD said unto Moses, Stretch out thine

hand toward heaven, that there may be darkness over the land of Egypt, even darkness which may be felt.

num tuam in caelum et sint tenebrae super terram Aegypti tam densae ut palpari queant 10:22 extendit Moses manum 10:22 And Moses stretched forth his hand toward heaven;

and there was a thick darkness in all the land of Egypt three days:

in caelum et factae sunt tenebrae horribiles in uniuersa terra Aegypti tribus diebus 10:23 nemo uidit fratrem suum nec 10:23 They saw not one another, neither rose any from his

place for three days: but all the children of Israel had light in their dwellings.

mouit se de loco in quo erat ubicumque autem habitabant filii Israhel lux erat 10:24 uocauitque Pharao Mosen et Aa- 10:24 And Pharaoh called unto Moses, and said, Go ye,

serve the LORD; only let your flocks and your herds be stayed: let your little ones also go with you.

ron et dixit eis ite sacrificate Domino oues tantum uestrae et armenta remaneant paruuli uestri eant uobiscum 10:25 ait 10:25 And Moses said, Thou must give us also sacrifices

and burnt offerings, that we may sacrifice unto the LORD our God.

Moses hostias quoque et holocausta dabis nobis quae offeramus Domino Deo nostro 10:26 cuncti greges pergent no- 10:26 Our cattle also shall go with us; there shall not an

hoof be left behind; for thereof must we take to serve the LORD our God; and we know not with what we must

biscum non remanebit ex eis ungula quae necessaria sunt in serve the LORD, until we come thither. cultum Domini Dei nostri praesertim cum ignoremus quid debeat immolari donec ad ipsum locum perueniamus 10:27 10:27 But the LORD hardened Pharaoh’s heart, and he

would not let them go.

indurauit autem Dominus cor Pharaonis et noluit dimittere

Biblia Sacra Vulgata

10:28 And Pharaoh said unto him, Get thee from me, take eos 10:28 dixitque Pharao ad eum recede a me caue ne ultra

heed to thyself, see my face no more; for in that day thou seest my face thou shalt die.

uideas faciem meam quocumque die apparueris mihi morie-

10:29 And Moses said, Thou hast spoken well, I will see ris 10:29 respondit Moses ita fiat ut locutus es non uidebo

thy face again no more.

ultra faciem tuam

11:1 And the LORD said unto Moses, Yet will I bring one 11:1 et dixit Dominus ad Mosen adhuc una plaga tangam plague more upon Pharaoh, and upon Egypt; afterwards he will let you go hence: when he shall let you go, he shall

surely thrust you out hence altogether. Pharaonem et Aegyptum et post haec dimittet uos et exire 11:2 Speak now in the ears of the people, and let every conpellet 11:2 dices ergo omni plebi ut postulet uir ab am-

man borrow of his neighbour, and every woman of her neighbour, jewels of silver and jewels of gold.

ico suo et mulier a uicina sua uasa argentea et aurea 11:3

11:3 And the LORD gave the people favour in the sight of the Egyptians. Moreover the man Moses was very great in the land of Egypt, in the sight of Pharaoh’s servants, and in the sight of the people.

dabit autem Dominus gratiam populo coram Aegyptiis fuitque Moses uir magnus ualde in terra Aegypti coram seruis

11:4 And Moses said, Thus saith the LORD, About mid- Pharao et omni populo 11:4 et ait haec dicit Dominus me-

night will I go out into the midst of Egypt:

11:5 And all the firstborn in the land of Egypt shall die, dia nocte egrediar in Aegyptum 11:5 et morietur omne pri-

from the first born of Pharaoh that sitteth upon his throne, even unto the firstborn of the maidservant that is behind

the mill; and all the firstborn of beasts. mogenitum in terra Aegyptiorum a primogenito Pharaonis

qui sedet in solio eius usque ad primogenitum ancillae quae

11:6 And there shall be a great cry throughout all the land est ad molam et omnia primogenita iumentorum 11:6 eritof Egypt, such as there was none like it, nor shall be like it any more.

que clamor magnus in uniuersa terra Aegypti qualis nec ante

11:7 But against any of the children of Israel shall not a fuit nec postea futurus est 11:7 apud omnes autem filios Is-

dog move his tongue, against man or beast: that ye may know how that the LORD doth put a difference between

the Egyptians and Israel. rahel non muttiet canis ab homine usque ad pecus ut sciatis 11:8 And all these thy servants shall come down unto me, quanto miraculo diuidat Dominus Aegyptios et Israhel 11:8

and bow down themselves unto me, saying, Get thee out, and all the people that follow thee: and after that I will go

out. And he went out from Pharaoh in a great anger. descendentque omnes serui tui isti ad me et adorabunt me

dicentes egredere tu et omnis populus qui subiectus est tibi

11:9 And the LORD said unto Moses, Pharaoh shall not post haec egrediemur 11:9 et exiuit a Pharaone iratus ni-

hearken unto you; that my wonders may be multiplied in the land of Egypt.

mis dixit autem Dominus ad Mosen non audiet uos Pharao

11:10 And Moses and Aaron did all these wonders before ut multa signa fiant in terra Aegypti 11:10 Moses autem et

Pharaoh: and the LORD hardened Pharaoh’s heart, so that he would not let the children of Israel go out of his land.

Aaron fecerunt omnia ostenta quae scripta sunt coram Pharaone et indurauit Dominus cor Pharaonis nec dimisit filios Israhel de terra sua

12:1 And the LORD spake unto Moses and Aaron in the 12:1 dixit quoque Dominus ad Mosen et Aaron in terra Ae-

land of Egypt saying,

12:2 This month shall be unto you the beginning of gypti 12:2 mensis iste uobis principium mensuum primus

months: it shall be the first month of the year to you. 12:3 Speak ye unto all the congregation of Israel, saying, erit in mensibus anni 12:3 loquimini ad uniuersum coetum In the tenth day of this month they shall take to them every man a lamb, according to the house of their fathers, a lamb

for an house: filiorum Israhel et dicite eis decima die mensis huius tol-

12:4 And if the household be too little for the lamb, let him lat unusquisque agnum per familias et domos suas 12:4 sin and his neighbour next unto his house take it according to the number of the souls; every man according to his eating

shall make your count for the lamb. autem minor est numerus ut sufficere possit ad uescendum agnum adsumet uicinum suum qui iunctus est domui eius iuxta numerum animarum quae sufficere possunt ad esum agni 12:5 erit autem agnus absque macula masculus anni- 12:5 Your lamb shall be without blemish, a male of the

first year: ye shall take it out from the sheep, or from the goats:

culus iuxta quem ritum tolletis et hedum 12:6 et seruabitis

12:6 And ye shall keep it up until the fourteenth day of the same month: and the whole assembly of the congregation of Israel shall kill it in the evening.

eum usque ad quartamdecimam diem mensis huius immolabitque eum uniuersa multitudo filiorum Israhel ad uesperam 12:7 et sument de sanguine ac ponent super utrumque postem 12:7 And they shall take of the blood, and strike it on the

two side posts and on the upper door post of the houses, wherein they shall eat it.

et in superliminaribus domorum in quibus comedent illum 12:8 et edent carnes nocte illa assas igni et azymos panes 12:8 And they shall eat the flesh in that night, roast with

fire, and unleavened bread; and with bitter herbs they shall eat it.

cum lactucis agrestibus 12:9 non comedetis ex eo crudum

12:9 Eat not of it raw, nor sodden at all with water, but roast with fire; his head with his legs, and with the purtenance thereof.

quid nec coctum aqua sed assum tantum igni caput cum pedibus eius et intestinis uorabitis 12:10 nec remanebit ex eo 12:10 And ye shall let nothing of it remain until the morn-

ing; and that which remaineth of it until the morning ye shall burn with fire.

quicquam usque mane si quid residui fuerit igne conburetis 12:11 sic autem comedetis illum renes uestros accingetis cal- 12:11 And thus shall ye eat it; with your loins girded, your

shoes on your feet, and your staff in your hand; and ye shall eat it in haste: it is the LORD’s passover.

ciamenta habebitis in pedibus tenentes baculos in manibus et comedetis festinantes est enim phase id est transitus Domini 12:12 et transibo per terram Aegypti nocte illa percutiamque 12:12 For I will pass through the land of Egypt this night,

and will smite all the firstborn in the land of Egypt, both man and beast; and against all the gods of Egypt I will

omne primogenitum in terra Aegypti ab homine usque ad execute judgment: I am the LORD. pecus et in cunctis diis Aegypti faciam iudicia ego Dominus 12:13 erit autem sanguis uobis in signum in aedibus in 12:13 And the blood shall be to you for a token upon the

houses where ye are: and when I see the blood, I will pass over you, and the plague shall not be upon you to destroy

quibus eritis et uidebo sanguinem ac transibo uos nec erit in you, when I smite the land of Egypt. uobis plaga disperdens quando percussero terram Aegypti 12:14 habebitis autem hanc diem in monumentum et cele- 12:14 And this day shall be unto you for a memorial; and

ye shall keep it a feast to the LORD throughout your generations; ye shall keep it a feast by an ordinance for ever.

brabitis eam sollemnem Domino in generationibus uestris cultu sempiterno 12:15 septem diebus azyma comedetis in 12:15 Seven days shall ye eat unleavened bread; even the

first day ye shall put away leaven out of your houses: for whosoever eateth leavened bread from the first day until

die primo non erit fermentum in domibus uestris quicumque the seventh day, that soul shall be cut off from Israel. comederit fermentatum peribit anima illa de Israhel a primo die usque ad diem septimum 12:16 dies prima erit sancta 12:16 And in the first day there shall be an holy convoca-

tion, and in the seventh day there shall be an holy convocation to you; no manner of work shall be done in them, save that which every man must eat, that only may be done of you.

atque sollemnis et dies septima eadem festiuitate uenerabilis nihil operis facietis in eis exceptis his quae ad uescendum pertinent 12:17 et obseruabitis azyma in eadem enim ipsa 12:17 And ye shall observe the feast of unleavened bread;

for in this selfsame day have I brought your armies out of the land of Egypt: therefore shall ye observe this day in

die educam exercitum uestrum de terra Aegypti et custodie- your generations by an ordinance for ever. tis diem istum in generationes uestras ritu perpetuo 12:18 12:18 In the first month, on the fourteenth day of the

month at even, ye shall eat unleavened bread, until the one and twentieth day of the month at even.

Biblia Sacra Vulgata

primo mense quartadecima die mensis ad uesperam comedetis azyma usque ad diem uicesimam primam eiusdem mensis

12:19 Seven days shall there be no leaven found in your ad uesperam 12:19 septem diebus fermentum non inuenie-

houses: for whosoever eateth that which is leavened, even that soul shall be cut off from the congregation of Israel,

whether he be a stranger, or born in the land. tur in domibus uestris qui comederit fermentatum peribit an-

ima eius de coetu Israhel tam de aduenis quam de indigenis

12:20 Ye shall eat nothing leavened; in all your habitations terrae 12:20 omne fermentatum non comedetis in cunctis shall ye eat unleavened bread.

12:21 Then Moses called for all the elders of Israel, and habitaculis uestris edetis azyma 12:21 uocauit autem Moses

said unto them, Draw out and take you a lamb according to your families, and kill the passover.

omnes seniores filiorum Israhel et dixit ad eos ite tollentes

12:22 And ye shall take a bunch of hyssop, and dip it in the animal per familias uestras immolate phase 12:22 fascicu-

blood that is in the bason, and strike the lintel and the two side posts with the blood that is in the bason; and none of

you shall go out at the door of his house until the morning. lumque hysopi tinguite sanguine qui est in limine et asper-

gite ex eo superliminare et utrumque postem nullus uestrum

12:23 For the LORD will pass through to smite the Egyp- egrediatur ostium domus suae usque mane 12:23 transibit

tians; and when he seeth the blood upon the lintel, and on the two side posts, the LORD will pass over the door, and will not suffer the destroyer to come in unto your houses to smite you.

enim Dominus percutiens Aegyptios cumque uiderit sanguinem in superliminari et in utroque poste transcendet ostium et non sinet percussorem ingredi domos uestras et laedere

12:24 And ye shall observe this thing for an ordinance to 12:24 custodi uerbum istud legitimum tibi et filiis tuis usque

thee and to thy sons for ever.

12:25 And it shall come to pass, when ye be come to the in aeternum 12:25 cumque introieritis terram quam Domi-

land which the LORD will give you, according as he hath promised, that ye shall keep this service.

nus daturus est uobis ut pollicitus est obseruabitis caerimo-

12:26 And it shall come to pass, when your children shall nias istas 12:26 et cum dixerint uobis filii uestri quae est

say unto you, What mean ye by this service?

12:27 That ye shall say, It is the sacrifice of the LORD’s ista religio 12:27 dicetis eis uictima transitus Domini est

passover, who passed over the houses of the children of Israel in Egypt, when he smote the Egyptians, and delivered our houses. And the people bowed the head and worshipped.

quando transiuit super domos filiorum Israhel in Aegypto percutiens Aegyptios et domos nostras liberans incuruatus-

12:28 And the children of Israel went away, and did as the que populus adorauit 12:28 et egressi filii Israhel fecerunt

LORD had commanded Moses and Aaron, so did they.

12:29 And it came to pass, that at midnight the LORD sicut praeceperat Dominus Mosi et Aaron 12:29 factum est

smote all the firstborn in the land of Egypt, from the firstborn of Pharaoh that sat on his throne unto the firstborn of the captive that was in the dungeon; and all the firstborn of cattle.

autem in noctis medio percussit Dominus omne primogenitum in terra Aegypti a primogenito Pharaonis qui sedebat in solio eius usque ad primogenitum captiuae quae erat in car-

12:30 And Pharaoh rose up in the night, he, and all his cere et omne primogenitum iumentorum 12:30 surrexitque

servants, and all the Egyptians; and there was a great cry in Egypt; for there was not a house where there was not

one dead. Pharao nocte et omnes serui eius cunctaque Aegyptus et or-

tus est clamor magnus in Aegypto neque enim erat domus in

12:31 And he called for Moses and Aaron by night, and qua non iaceret mortuus 12:31 uocatisque Mosen et Aaron

said, Rise up, and get you forth from among my people, both ye and the children of Israel; and go, serve the LORD,

as ye have said. nocte ait surgite egredimini a populo meo et uos et filii Is-12:32 Also take your flocks and your herds, as ye have rahel ite immolate Domino sicut dicitis 12:32 oues uestras

said, and be gone; and bless me also.

et armenta adsumite ut petieratis et abeuntes benedicite mihi 12:33 urguebantque Aegyptii populum de terra exire ueloci- 12:33 And the Egyptians were urgent upon the people, that

they might send them out of the land in haste; for they said, We be all dead men.

ter dicentes omnes moriemur 12:34 tulit igitur populus con-

12:34 And the people took their dough before it was leavened, their kneadingtroughs being bound up in their clothes upon their shoulders.

spersam farinam antequam fermentaretur et ligans in palliis posuit super umeros suos 12:35 feceruntque filii Israhel sic- 12:35 And the children of Israel did according to the word

of Moses; and they borrowed of the Egyptians jewels of silver, and jewels of gold, and raiment:

ut praeceperat Moses et petierunt ab Aegyptiis uasa argentea et aurea uestemque plurimam 12:36 dedit autem Dominus 12:36 And the LORD gave the people favour in the sight

of the Egyptians, so that they lent unto them such things as they required. And they spoiled the Egyptians.

gratiam populo coram Aegyptiis ut commodarent eis et spoliauerunt Aegyptios 12:37 profectique sunt filii Israhel de 12:37 And the children of Israel journeyed from Rameses

to Succoth, about six hundred thousand on foot that were men, beside children.

Ramesse in Soccoth sescenta ferme milia peditum uirorum absque paruulis 12:38 sed et uulgus promiscuum innume- 12:38 And a mixed multitude went up also with them; and

flocks, and herds, even very much cattle.

rabile ascendit cum eis oues et armenta et animantia diuersi generis multa nimis 12:39 coxeruntque farinam quam du- 12:39 And they baked unleavened cakes of the dough

which they brought forth out of Egypt, for it was not leavened; because they were thrust out of Egypt, and could not

dum conspersam de Aegypto tulerant et fecerunt subcineri- tarry, neither had they prepared for themselves any victual.

cios panes azymos neque enim poterant fermentari cogentibus exire Aegyptiis et nullam facere sinentibus moram nec pulmenti quicquam occurrerant praeparare 12:40 habitatio 12:40 Now the sojourning of the children of Israel, who

dwelt in Egypt, was four hundred and thirty years.

autem filiorum Israhel qua manserant in Aegypto fuit quadringentorum triginta annorum 12:41 quibus expletis eadem 12:41 And it came to pass at the end of the four hundred

and thirty years, even the selfsame day it came to pass, that all the hosts of the LORD went out from the land of

die egressus est omnis exercitus Domini de terra Aegypti Egypt. 12:42 nox est ista obseruabilis Domini quando eduxit eos de 12:42 It is a night to be much observed unto the LORD for

bringing them out from the land of Egypt: this is that night of the LORD to be observed of all the children of Israel in

terra Aegypti hanc obseruare debent omnes filii Israhel in their generations. generationibus suis 12:43 dixitque Dominus ad Mosen et 12:43 And the LORD said unto Moses and Aaron, This is

the ordinance of the passover: There shall no stranger eat thereof:

Aaron haec est religio phase omnis alienigena non comedet ex eo 12:44 omnis autem seruus empticius circumcidetur 12:44 But every man’s servant that is bought for money,

when thou hast circumcised him, then shall he eat thereof.

et sic comedet 12:45 aduena et mercennarius non edent ex 12:45 A foreigner and an hired servant shall not eat

thereof.

eo 12:46 in una domo comedetur nec efferetis de carnibus 12:46 In one house shall it be eaten; thou shalt not carry

forth ought of the flesh abroad out of the house; neither shall ye break a bone thereof.

eius foras nec os illius confringetis 12:47 omnis coetus fi-

12:47 All the congregation of Israel shall keep it.

liorum Israhel faciet illud 12:48 quod si quis peregrinorum 12:48 And when a stranger shall sojourn with thee, and

will keep the passover to the LORD, let all his males be circumcised, and then let him come near and keep it; and he shall be as one that is born in the land: for no uncircumcised person shall eat thereof.

in uestram uoluerit transire coloniam et facere phase Domini circumcidetur prius omne masculinum eius et tunc rite celebrabit eritque sicut indigena terrae si quis autem circumcisus non fuerit non uescetur ex eo 12:49 eadem lex erit indige- 12:49 One law shall be to him that is homeborn, and unto

the stranger that sojourneth among you.

Biblia Sacra Vulgata

12:50 Thus did all the children of Israel; as the LORD nae et colono qui peregrinatur apud uos 12:50 fecerunt om-

commanded Moses and Aaron, so did they.

nes filii Israhel sicut praeceperat Dominus Mosi et Aaron

12:51 And it came to pass the selfsame day, that the LORD 12:51 et in eadem die eduxit Dominus filios Israhel de terra

did bring the children of Israel out of the land of Egypt by their armies.

Aegypti per turmas suas

13:1 And the LORD spake unto Moses, saying, 13:1 locutusque est Dominus ad Mosen dicens 13:2 sanc-

13:2 Sanctify unto me all the firstborn, whatsoever openeth the womb among the children of Israel, both of man and of beast: it is mine.

tifica mihi omne primogenitum quod aperit uuluam in filiis Israhel tam de hominibus quam de iumentis mea sunt

13:3 And Moses said unto the people, Remember this day, enim omnia 13:3 et ait Moses ad populum mementote diei in which ye came out from Egypt, out of the house of bondage; for by strength of hand the LORD brought you out from this place: there shall no leavened bread be eaten. huius in qua egressi estis de Aegypto et de domo seruitutis

quoniam in manu forti eduxit uos Dominus de loco isto ut

13:4 This day came ye out in the month Abib. non comedatis fermentatum panem 13:4 hodie egredimini

13:5 And it shall be when the LORD shall bring thee into mense nouarum frugum 13:5 cumque te introduxerit Dothe land of the Canaanites, and the Hittites, and the Amorites, and the Hivites, and the Jebusites, which he sware unto thy fathers to give thee, a land flowing with milk and honey, that thou shalt keep this service in this month.

minus in terram Chananei et Hetthei et Amorrei et Euei et Iebusei quam iurauit patribus tuis ut daret tibi terram fluentem lacte et melle celebrabis hunc morem sacrorum mense

13:6 Seven days thou shalt eat unleavened bread, and in isto 13:6 septem diebus uesceris azymis et in die septimo

the seventh day shall be a feast to the LORD.

13:7 Unleavened bread shall be eaten seven days; and erit sollemnitas Domini 13:7 azyma comedetis septem die-

there shall no leavened bread be seen with thee, neither shall there be leaven seen with thee in all thy quarters.

bus non apparebit apud te aliquid fermentatum nec in cunctis

13:8 And thou shalt shew thy son in that day, saying, This finibus tuis 13:8 narrabisque filio tuo in die illo dicens hoc

is done because of that which the LORD did unto me when I came forth out of Egypt.

est quod fecit Dominus mihi quando egressus sum de Ae-

13:9 And it shall be for a sign unto thee upon thine hand, gypto 13:9 et erit quasi signum in manu tua et quasi monu-

and for a memorial between thine eyes, that the LORD’s law may be in thy mouth: for with a strong hand hath the

LORD brought thee out of Egypt. mentum ante oculos tuos et ut lex Domini semper in ore tuo

13:10 Thou shalt therefore keep this ordinance in his sea- in manu enim forti eduxit te Dominus de Aegypto 13:10 cuson from year to year.

stodies huiuscemodi cultum statuto tempore a diebus in dies

13:11 And it shall be when the LORD shall bring thee into 13:11 cumque introduxerit te in terram Chananei sicut iura-

the land of the Canaanites, as he sware unto thee and to thy fathers, and shall give it thee,

uit tibi et patribus tuis et dederit eam tibi 13:12 separabis

13:12 That thou shalt set apart unto the LORD all that openeth the matrix, and every firstling that cometh of a beast which thou hast; the males shall be the LORD’s.

omne quod aperit uuluam Domino et quod primitiuum est in pecoribus tuis quicquid habueris masculini sexus consecra-

13:13 And every firstling of an ass thou shalt redeem with bis Domino 13:13 primogenitum asini mutabis oue quod si a lamb; and if thou wilt not redeem it, then thou shalt break his neck: and all the firstborn of man among thy children

shalt thou redeem. non redeeeris interficies omne autem primogenitum hominis 13:14 And it shall be when thy son asketh thee in time to de filiis tuis pretio rediies 13:14 cumque interrogauerit te

come, saying, What is this? that thou shalt say unto him, By strength of hand the LORD brought us out from Egypt,

from the house of bondage: filius tuus cras dicens quid est hoc respondebis ei in manu forti eduxit nos Dominus de Aegypto de domo seruitutis 13:15 nam cum induratus esset Pharao et nollet nos dimit- 13:15 And it came to pass, when Pharaoh would hardly

let us go, that the LORD slew all the firstborn in the land of Egypt, both the firstborn of man, and the firstborn of beast: therefore I sacrifice to the LORD all that openeth the matrix, being males; but all the firstborn of my children I redeem.

tere occidit Dominus omne primogenitum in terra Aegypti a primogenito hominis usque ad primogenitum iumentorum idcirco immolo Domino omne quod aperit uuluam masculini sexus et omnia primogenita filiorum meorum rediio 13:16 13:16 And it shall be for a token upon thine hand, and for

frontlets between thine eyes: for by strength of hand the LORD brought us forth out of Egypt.

erit igitur quasi signum in manu tua et quasi adpensum quid ob recordationem inter oculos tuos eo quod in manu forti eduxerit nos Dominus de Aegypto 13:17 igitur cum emisis- 13:17 And it came to pass, when Pharaoh had let the peo-

ple go, that God led them not through the way of the land of the Philistines, although that was near; for God said, Lest peradventure the people repent when they see war, and they return to Egypt:

set Pharao populum non eos duxit Dominus per uiam terrae Philisthim quae uicina est reputans ne forte paeniteret eum si uidisset aduersum se bella consurgere et reuerteretur in Aegyptum 13:18 sed circumduxit per uiam deserti quae 13:18 But God led the people about, through the way of

the wilderness of the Red sea: and the children of Israel went up harnessed out of the land of Egypt.

est iuxta mare Rubrum et armati ascenderunt filii Israhel de terra Aegypti 13:19 tulit quoque Moses ossa Ioseph secum 13:19 And Moses took the bones of Joseph with him: for

he had straitly sworn the children of Israel, saying, God will surely visit you; and ye shall carry up my bones away

eo quod adiurasset filios Israhel dicens uisitabit uos Deus hence with you. efferte ossa mea hinc uobiscum 13:20 profectique de Soc- 13:20 And they took their journey from Succoth, and en-

camped in Etham, in the edge of the wilderness.

coth castrametati sunt in Etham in extremis finibus solitudinis 13:21 Dominus autem praecedebat eos ad ostendendam 13:21 And the LORD went before them by day in a pillar

of a cloud, to lead them the way; and by night in a pillar of fire, to give them light; to go by day and night:

uiam per diem in columna nubis et per noctem in columna ignis ut dux esset itineris utroque tempore 13:22 numquam 13:22 He took not away the pillar of the cloud by day, nor

the pillar of fire by night, from before the people.

defuit columna nubis per diem nec columna ignis per noctem coram populo

14:1 locutus est autem Dominus ad Mosen dicens 14:2 14:1 And the LORD spake unto Moses, saying,

14:2 Speak unto the children of Israel, that they turn and encamp before Pihahiroth, between Migdol and the sea, over against Baalzephon: before it shall ye encamp by the sea.

loquere filiis Israhel reuersi castrametentur e regione Phiahiroth quae est inter Magdolum et mare contra Beelsephon in conspectu eius castra ponetis super mare 14:3 dicturusque 14:3 For Pharaoh will say of the children of Israel, They

are entangled in the land, the wilderness hath shut them in.

est Pharao super filiis Israhel coartati sunt in terra conclusit eos desertum 14:4 et indurabo cor eius ac persequetur 14:4 And I will harden Pharaoh’s heart, that he shall fol-

low after them; and I will be honoured upon Pharaoh, and upon all his host; that the Egyptians may know that I am

uos et glorificabor in Pharao et in omni exercitu eius scient- the LORD. And they did so.

que Aegyptii quia ego sum Dominus feceruntque ita 14:5 14:5 And it was told the king of Egypt that the people fled: and the heart of Pharaoh and of his servants was turned against the people, and they said, Why have we done this,

et nuntiatum est regi Aegyptiorum quod fugisset populus in- that we have let Israel go from serving us? mutatumque est cor Pharaonis et seruorum eius super populo et dixerunt quid uoluimus facere ut dimitteremus Israhel ne

Biblia Sacra Vulgata

14:6 And he made ready his chariot, and took his people seruiret nobis 14:6 iunxit ergo currum et omnem populum

with him:

14:7 And he took six hundred chosen chariots, and all the suum adsumpsit secum 14:7 tulitque sescentos currus elec-

chariots of Egypt, and captains over every one of them.

tos quicquid in Aegypto curruum fuit et duces totius exer-

14:8 And the LORD hardened the heart of Pharaoh king citus 14:8 indurauitque Dominus cor Pharaonis regis Ae-

of Egypt, and he pursued after the children of Israel: and the children of Israel went out with an high hand.

gypti et persecutus est filios Israhel at illi egressi erant in

14:9 But the Egyptians pursued after them, all the horses manu excelsa 14:9 cumque persequerentur Aegyptii uesti-

and chariots of Pharaoh, and his horsemen, and his army, and overtook them encamping by the sea, beside Pihahi-

roth, before Baalzephon. gia praecedentium reppererunt eos in castris super mare om-

nis equitatus et currus Pharaonis et uniuersus exercitus er-

14:10 And when Pharaoh drew nigh, the children of Israel ant in Ahiroth contra Beelsephon 14:10 cumque adpropin-

lifted up their eyes, and, behold, the Egyptians marched after them; and they were sore afraid: and the children of

Israel cried out unto the LORD. quasset Pharao leuantes filii Israhel oculos uiderunt Aegyp-

tios post se et timuerunt ualde clamaueruntque ad Dominum

14:11 And they said unto Moses, Because there were no 14:11 et dixerunt ad Mosen forsitan non erant sepulchra in

graves in Egypt, hast thou taken us away to die in the wilderness? wherefore hast thou dealt thus with us, to

carry us forth out of Egypt? Aegypto ideo tulisti nos ut moreremur in solitudine quid hoc 14:12 Is not this the word that we did tell thee in Egypt, facere uoluisti ut educeres nos ex Aegypto 14:12 nonne iste

saying, Let us alone, that we may serve the Egyptians? For it had been better for us to serve the Egyptians, than that

we should die in the wilderness. est sermo quem loquebamur ad te in Aegypto dicentes re-

cede a nobis ut seruiamus Aegyptiis multo enim melius est

14:13 And Moses said unto the people, Fear ye not, stand seruire eis quam mori in solitudine 14:13 et ait Moses ad

still, and see the salvation of the LORD, which he will shew to you to day: for the Egyptians whom ye have seen

to day, ye shall see them again no more for ever. populum nolite timere state et uidete magnalia Domini quae

facturus est hodie Aegyptios enim quos nunc uidetis nequa-

14:14 The LORD shall fight for you, and ye shall hold quam ultra uidebitis usque in sempiternum 14:14 Dominus

your peace.

14:15 And the LORD said unto Moses, Wherefore criest pugnabit pro uobis et uos tacebitis 14:15 dixitque Domi-

thou unto me? speak unto the children of Israel, that they go forward:

nus ad Mosen quid clamas ad me loquere filiis Israhel ut

14:16 But lift thou up thy rod, and stretch out thine hand proficiscantur 14:16 tu autem eleua uirgam tuam et extende

over the sea, and divide it: and the children of Israel shall go on dry ground through the midst of the sea.

manum super mare et diuide illud ut gradiantur filii Israhel

14:17 And I, behold, I will harden the hearts of the Egyp- in medio mari per siccum 14:17 ego autem indurabo cor tians, and they shall follow them: and I will get me honour upon Pharaoh, and upon all his host, upon his chariots, and

upon his horsemen. Aegyptiorum ut persequantur uos et glorificabor in Phara-

one et in omni exercitu eius in curribus et in equitibus illius

14:18 And the Egyptians shall know that I am the LORD, 14:18 et scient Aegyptii quia ego sum Dominus cum glorifi-

when I have gotten me honour upon Pharaoh, upon his chariots, and upon his horsemen.

catus fuero in Pharaone et in curribus atque in equitibus eius

14:19 And the angel of God, which went before the camp 14:19 tollensque se angelus Dei qui praecedebat castra Isra-

of Israel, removed and went behind them; and the pillar of the cloud went from before their face, and stood behind

them: hel abiit post eos et cum eo pariter columna nubis priora di-14:20 And it came between the camp of the Egyptians and mittens post tergum 14:20 stetit inter castra Aegyptiorum et

the camp of Israel; and it was a cloud and darkness to them, but it gave light by night to these: so that the one

came not near the other all the night. castra Israhel et erat nubes tenebrosa et inluminans noctem ut ad se inuicem toto noctis tempore accedere non ualerent 14:21 cumque extendisset Moses manum super mare abstu- 14:21 And Moses stretched out his hand over the sea; and

the LORD caused the sea to go back by a strong east wind all that night, and made the sea dry land, and the waters

lit illud Dominus flante uento uehementi et urente tota nocte were divided. et uertit in siccum diuisaque est aqua 14:22 et ingressi sunt 14:22 And the children of Israel went into the midst of the

sea upon the dry ground: and the waters were a wall unto them on their right hand, and on their left.

filii Israhel per medium maris sicci erat enim aqua quasi murus a dextra eorum et leua 14:23 persequentesque Aegyptii 14:23 And the Egyptians pursued, and went in after them

to the midst of the sea, even all Pharaoh’s horses, his chariots, and his horsemen.

ingressi sunt post eos omnis equitatus Pharaonis currus eius et equites per medium maris 14:24 iamque aduenerat uigilia 14:24 And it came to pass, that in the morning watch the

LORD looked unto the host of the Egyptians through the pillar of fire and of the cloud, and troubled the host of the

matutina et ecce respiciens Dominus super castra Aegyptio- Egyptians, rum per columnam ignis et nubis interfecit exercitum eorum 14:25 et subuertit rotas curruum ferebanturque in profundum 14:25 And took off their chariot wheels, that they drave

them heavily: so that the Egyptians said, Let us flee from the face of Israel; for the LORD fighteth for them against

dixerunt ergo Aegyptii fugiamus Israhelem Dominus enim the Egyptians. pugnat pro eis contra nos 14:26 et ait Dominus ad Mosen 14:26 And the LORD said unto Moses, Stretch out thine

hand over the sea, that the waters may come again upon the Egyptians, upon their chariots, and upon their horse-

extende manum tuam super mare ut reuertantur aquae ad Ae- men. gyptios super currus et equites eorum 14:27 cumque exten- 14:27 And Moses stretched forth his hand over the sea,

and the sea returned to his strength when the morning appeared; and the Egyptians fled against it; and the LORD

disset Moses manum contra mare reuersum est primo dilu- overthrew the Egyptians in the midst of the sea. culo ad priorem locum fugientibusque Aegyptiis occurrerunt aquae et inuoluit eos Dominus in mediis fluctibus 14:28 re- 14:28 And the waters returned, and covered the chariots,

and the horsemen, and all the host of Pharaoh that came into the sea after them; there remained not so much as one

uersaeque sunt aquae et operuerunt currus et equites cuncti of them. exercitus Pharaonis qui sequentes ingressi fuerant mare ne unus quidem superfuit ex eis 14:29 filii autem Israhel per- 14:29 But the children of Israel walked upon dry land in

the midst of the sea; and the waters were a wall unto them on their right hand, and on their left.

rexerunt per medium sicci maris et aquae eis erant quasi pro muro a dextris et a sinistris 14:30 liberauitque Dominus in 14:30 Thus the LORD saved Israel that day out of the hand

of the Egyptians; and Israel saw the Egyptians dead upon the sea shore.

die illo Israhel de manu Aegyptiorum 14:31 et uiderunt Ae-

14:31 And Israel saw that great work which the LORD did upon the Egyptians: and the people feared the LORD, and believed the LORD, and his servant Moses.

gyptios mortuos super litus maris et manum magnam quam exercuerat Dominus contra eos timuitque populus Dominum et crediderunt Domino et Mosi seruo eius

15:1 tunc cecinit Moses et filii Israhel carmen hoc Do- 15:1 Then sang Moses and the children of Israel this song unto the LORD, and spake, saying, I will sing unto the LORD, for he hath triumphed gloriously: the horse and

mino et dixerunt cantemus Domino gloriose enim magni- his rider hath he thrown into the sea. ficatus est equum et ascensorem deiecit in mare 15:2 forti- 15:2 The LORD is my strength and song, and he is become

my salvation: he is my God, and I will prepare him an habitation; my father’s God, and I will exalt him.

tudo mea et laus mea Dominus et factus est mihi in salutem iste Deus meus et glorificabo eum Deus patris mei et exaltabo eum 15:3 Dominus quasi uir pugnator Omnipotens 15:3 The LORD is a man of war: the LORD is his name.

Biblia Sacra Vulgata

15:4 Pharaoh’s chariots and his host hath he cast into the nomen eius 15:4 currus Pharaonis et exercitum eius proie-

sea: his chosen captains also are drowned in the Red sea.

cit in mare electi principes eius submersi sunt in mari Rubro

15:5 The depths have covered them: they sank into the 15:5 abyssi operuerunt eos descenderunt in profundum quasi

bottom as a stone.

15:6 Thy right hand, O LORD, is become glorious in lapis 15:6 dextera tua Domine magnifice in fortitudine de-

power: thy right hand, O LORD, hath dashed in pieces the enemy.

xtera tua Domine percussit inimicum 15:7 et in multitudine

15:7 And in the greatness of thine excellency thou hast overthrown them that rose up against thee: thou sentest forth thy wrath, which consumed them as stubble.

gloriae tuae deposuisti aduersarios meos misisti iram tuam

15:8 And with the blast of thy nostrils the waters were quae deuorauit eos ut stipulam 15:8 et in spiritu furoris tui

gathered together, the floods stood upright as an heap, and the depths were congealed in the heart of the sea.

congregatae sunt aquae stetit unda fluens congregatae sunt

15:9 The enemy said, I will pursue, I will overtake, I will abyssi in medio mari 15:9 dixit inimicus persequar et con-

divide the spoil; my lust shall be satisfied upon them; I will draw my sword, my hand shall destroy them.

prehendam diuidam spolia implebitur anima mea euaginabo

15:10 Thou didst blow with thy wind, the sea covered gladium meum interficiet eos manus mea 15:10 flauit spi-

them: they sank as lead in the mighty waters.

ritus tuus et operuit eos mare submersi sunt quasi plumbum

15:11 Who is like unto thee, O LORD, among the gods? in aquis uehementibus 15:11 quis similis tui in fortibus Do-

who is like thee, glorious in holiness, fearful in praises, doing wonders?

mine quis similis tui magnificus in sanctitate terribilis at-

15:12 Thou stretchedst out thy right hand, the earth swal- que laudabilis et faciens mirabilia 15:12 extendisti manum

lowed them.

15:13 Thou in thy mercy hast led forth the people which tuam et deuorauit eos terra 15:13 dux fuisti in misericordia

thou hast redeemed: thou hast guided them in thy strength unto thy holy habitation.

tua populo quem redeeisti et portasti eum in fortitudine tua

15:14 The people shall hear, and be afraid: sorrow shall ad habitaculum sanctum tuum 15:14 adtenderunt populi et

take hold on the inhabitants of Palestina.

15:15 Then the dukes of Edom shall be amazed; the irati sunt dolores obtinuerunt habitatores Philisthim 15:15

mighty men of Moab, trembling shall take hold upon them; all the inhabitants of Canaan shall melt away.

tunc conturbati sunt principes Edom robustos Moab obti-

15:16 Fear and dread shall fall upon them; by the greatness nuit tremor obriguerunt omnes habitatores Chanaan 15:16 of thine arm they shall be as still as a stone; till thy people pass over, O LORD, till the people pass over, which thou

hast purchased. inruat super eos formido et pauor in magnitudine brachii tui fiant inmobiles quasi lapis donec pertranseat populus tuus Domine donec pertranseat populus tuus iste quem posse-

15:17 Thou shalt bring them in, and plant them in the disti 15:17 introduces eos et plantabis in monte heredita-

mountain of thine inheritance, in the place, O LORD, which thou hast made for thee to dwell in, in the Sanc-

tuary, O LORD, which thy hands have established. tis tuae firmissimo habitaculo tuo quod operatus es Domine 15:18 The LORD shall reign for ever and ever. sanctuarium Domine quod firmauerunt manus tuae 15:18 15:19 For the horse of Pharaoh went in with his chariots Dominus regnabit in aeternum et ultra 15:19 ingressus est

and with his horsemen into the sea, and the LORD brought again the waters of the sea upon them; but the children of

Israel went on dry land in the midst of the sea. enim equus Pharao cum curribus et equitibus eius in mare

et reduxit super eos Dominus aquas maris filii autem Israhel

15:20 And Miriam the prophetess, the sister of Aaron, ambulauerunt per siccum in medio eius 15:20 sumpsit ergo

took a timbrel in her hand; and all the women went out after her with timbrels and with dances.

Maria prophetis soror Aaron tympanum in manu egressaeque sunt omnes mulieres post eam cum tympanis et choris 15:21 quibus praecinebat dicens cantemus Domino gloriose 15:21 And Miriam answered them, Sing ye to the LORD,

for he hath triumphed gloriously; the horse and his rider hath he thrown into the sea.

enim magnificatus est equum et ascensorem eius deiecit in mare 15:22 tulit autem Moses Israhel de mari Rubro et eg- 15:22 So Moses brought Israel from the Red sea, and they

went out into the wilderness of Shur; and they went three days in the wilderness, and found no water.

ressi sunt in desertum Sur ambulaueruntque tribus diebus per solitudinem et non inueniebant aquam 15:23 et uenerunt in 15:23 And when they came to Marah, they could not drink

of the waters of Marah, for they were bitter: therefore the name of it was called Marah.

Marath nec poterant bibere aquas de Mara eo quod essent amarae unde et congruum loco nomen inposuit uocans illud Mara id est amaritudinem 15:24 et murmurauit populus 15:24 And the people murmured against Moses, saying,

What shall we drink?

contra Mosen dicens quid bibemus 15:25 at ille clamauit 15:25 And he cried unto the LORD; and the LORD

shewed him a tree, which when he had cast into the waters, the waters were made sweet: there he made for them

ad Dominum qui ostendit ei lignum quod cum misisset in a statute and an ordinance, and there he proved them, aquas in dulcedinem uersae sunt ibi constituit ei praecepta atque iudicia et ibi temptauit eum 15:26 dicens si audieris 15:26 And said, If thou wilt diligently hearken to the voice

of the LORD thy God, and wilt do that which is right in his sight, and wilt give ear to his commandments, and keep all his statutes, I will put none of these diseases upon thee, which I have brought upon the Egyptians: for I am the LORD that healeth thee.

uocem Domini Dei tui et quod rectum est coram eo feceris et oboedieris mandatis eius custodierisque omnia praecepta illius cunctum languorem quem posui in Aegypto non inducam super te ego enim Dominus sanator tuus 15:27 uene- 15:27 And they came to Elim, where were twelve wells

of water, and threescore and ten palm trees: and they encamped there by the waters.

runt autem in Helim ubi erant duodecim fontes aquarum et septuaginta palmae et castrametati sunt iuxta aquas

16:1 profectique sunt de Helim et uenit omnis multitudo 16:1 And they took their journey from Elim, and all

the congregation of the children of Israel came unto the wilderness of Sin, which is between Elim and Sinai, on the fifteenth day of the second month after their departing out of the land of Egypt.

filiorum Israhel in desertum Sin quod est inter Helim et Sinai quintodecimo die mensis secundi postquam egressi sunt de terra Aegypti 16:2 et murmurauit omnis congregatio fi- 16:2 And the whole congregation of the children of Israel

murmured against Moses and Aaron in the wilderness:

liorum Israhel contra Mosen et contra Aaron in solitudine 16:3 dixeruntque ad eos filii Israhel utinam mortui essemus 16:3 And the children of Israel said unto them, Would to

God we had died by the hand of the LORD in the land of Egypt, when we sat by the flesh pots, and when we did eat bread to the full; for ye have brought us forth into this wilderness, to kill this whole assembly with hunger.

per manum Domini in terra Aegypti quando sedebamus super ollas carnium et comedebamus panes in saturitate cur eduxistis nos in desertum istud ut occideretis omnem multitudinem fame 16:4 dixit autem Dominus ad Mosen ecce 16:4 Then said the LORD unto Moses, Behold, I will rain

bread from heaven for you; and the people shall go out and gather a certain rate every day, that I may prove them,

ego pluam uobis panes de caelo egrediatur populus et colli- whether they will walk in my law, or no. gat quae sufficiunt per singulos dies ut temptem eum utrum

ambulet in lege mea an non 16:5 die autem sexta parent 16:5 And it shall come to pass, that on the sixth day they shall prepare that which they bring in; and it shall be twice as much as they gather daily.

quod inferant et sit duplum quam colligere solebant per singulos dies 16:6 dixeruntque Moses et Aaron ad omnes fi- 16:6 And Moses and Aaron said unto all the children of

Israel, At even, then ye shall know that the LORD hath brought you out from the land of Egypt:

Biblia Sacra Vulgata

lios Israhel uespere scietis quod Dominus eduxerit uos de

16:7 And in the morning, then ye shall see the glory of the terra Aegypti 16:7 et mane uidebitis gloriam Domini audi-

LORD; for that he heareth your murmurings against the LORD: and what are we, that ye murmur against us?

uit enim murmur uestrum contra Dominum nos uero quid

16:8 And Moses said, This shall be, when the LORD shall sumus quia mussitatis contra nos 16:8 et ait Moses dabit

give you in the evening flesh to eat, and in the morning bread to the full; for that the LORD heareth your murmurings which ye murmur against him: and what are we? your murmurings are not against us, but against the LORD.

Dominus uobis uespere carnes edere et mane panes in saturitate eo quod audierit murmurationes uestras quibus murmurati estis contra eum nos enim quid sumus nec contra

16:9 And Moses spake unto Aaron, Say unto all the con- nos est murmur uestrum sed contra Dominum 16:9 dixit-

gregation of the children of Israel, Come near before the LORD: for he hath heard your murmurings.

que Moses ad Aaron dic uniuersae congregationi filiorum Israhel accedite coram Domino audiuit enim murmur uestrum

16:10 And it came to pass, as Aaron spake unto the whole 16:10 cumque loqueretur Aaron ad omnem coetum filiorum

congregation of the children of Israel, that they looked toward the wilderness, and, behold, the glory of the LORD

appeared in the cloud. Israhel respexerunt ad solitudinem et ecce gloria Domini ap-16:11 And the LORD spake unto Moses, saying, paruit in nube 16:11 locutus est autem Dominus ad Mosen 16:12 I have heard the murmurings of the children of Is- dicens 16:12 audiui murmurationes filiorum Israhel loquere

rael: speak unto them, saying, At even ye shall eat flesh, and in the morning ye shall be filled with bread; and ye

shall know that I am the LORD your God. ad eos uespere comedetis carnes et mane saturabimini pani-16:13 And it came to pass, that at even the quails came bus scietisque quod sim Dominus Deus uester 16:13 factum

up, and covered the camp: and in the morning the dew lay round about the host.

est ergo uespere et ascendens coturnix operuit castra mane

16:14 And when the dew that lay was gone up, behold, quoque ros iacuit per circuitum castrorum 16:14 cumque

upon the face of the wilderness there lay a small round thing, as small as the hoar frost on the ground.

operuisset superficiem terrae apparuit in solitudine minutum et quasi pilo tunsum in similitudinem pruinae super terram

16:15 And when the children of Israel saw it, they said one 16:15 quod cum uidissent filii Israhel dixerunt ad inuicem

to another, It is manna: for they wist not what it was. And Moses said unto them, This is the bread which the LORD

hath given you to eat. man hu quod significat quid est hoc ignorabant enim quid

esset quibus ait Moses iste est panis quem dedit Dominus uo-

16:16 This is the thing which the LORD hath commanded, bis ad uescendum 16:16 hic est sermo quem praecepit Do-

Gather of it every man according to his eating, an omer for every man, according to the number of your persons; take

ye every man for them which are in his tents. minus colligat ex eo unusquisque quantum sufficiat ad ues-

cendum gomor per singula capita iuxta numerum animarum

16:17 And the children of Israel did so, and gathered, uestrarum quae habitant in tabernaculo sic tolletis 16:17 fe-

some more, some less.

ceruntque ita filii Israhel et collegerunt alius plus alius mi-

16:18 And when they did mete it with an omer, he that nus 16:18 et mensi sunt ad mensuram gomor nec qui plus

gathered much had nothing over, and he that gathered little had no lack; they gathered every man according to his

eating. collegerat habuit amplius nec qui minus parauerat repperit

minus sed singuli iuxta id quod edere poterant congregarunt

16:19 And Moses said, Let no man leave of it till the morn- 16:19 dixitque Moses ad eos nullus relinquat ex eo in mane

ing.

16:20 Notwithstanding they hearkened not unto Moses; 16:20 qui non audierunt eum sed dimiserunt quidam ex eis

but some of them left of it until the morning, and it bred worms, and stank: and Moses was wroth with them.

usque mane et scatere coepit uermibus atque conputruit et iratus est contra eos Moses 16:21 colligebant autem mane 16:21 And they gathered it every morning, every man ac-

cording to his eating: and when the sun waxed hot, it melted.

singuli quantum sufficere poterat ad uescendum cumque incaluisset sol liquefiebat 16:22 in die uero sexta collegerunt 16:22 And it came to pass, that on the sixth day they gath-

ered twice as much bread, two omers for one man: and all the rulers of the congregation came and told Moses.

cibos duplices id est duo gomor per singulos homines uenerunt autem omnes principes multitudinis et narrauerunt Mosi 16:23 qui ait eis hoc est quod locutus est Dominus requies 16:23 And he said unto them, This is that which the LORD

hath said, To morrow is the rest of the holy sabbath unto the LORD: bake that which ye will bake to day, and seethe that ye will seethe; and that which remaineth over lay up for you to be kept until the morning.

sabbati sanctificata erit Domino cras quodcumque operandum est facite et quae coquenda sunt coquite quicquid autem reliquum fuerit reponite usque in mane 16:24 fecerunt- 16:24 And they laid it up till the morning, as Moses bade:

and it did not stink, neither was there any worm therein.

que ita ut praeceperat Moses et non conputruit neque uermis inuentus est in eo 16:25 dixitque Moses comedite il- 16:25 And Moses said, Eat that to day; for to day is a

sabbath unto the LORD: to day ye shall not find it in the field.

lud hodie quia sabbatum est Domino non inuenietur hodie in agro 16:26 sex diebus colligite in die autem septimo sab- 16:26 Six days ye shall gather it; but on the seventh day,

which is the sabbath, in it there shall be none.

batum est Domino idcirco non inuenietur 16:27 uenit sep- 16:27 And it came to pass, that there went out some of the

people on the seventh day for to gather, and they found none.

tima dies et egressi de populo ut colligerent non inuenerunt 16:28 dixit autem Dominus ad Mosen usquequo non uultis 16:28 And the LORD said unto Moses, How long refuse

ye to keep my commandments and my laws?

custodire mandata mea et legem meam 16:29 uidete quod 16:29 See, for that the LORD hath given you the sabbath, therefore he giveth you on the sixth day the bread of two days; abide ye every man in his place, let no man go out

Dominus dederit uobis sabbatum et propter hoc tribuerit uo- of his place on the seventh day. bis die sexto cibos duplices maneat unusquisque apud semet ipsum nullus egrediatur de loco suo die septimo 16:30 16:30 So the people rested on the seventh day. et sabbatizauit populus die septimo 16:31 appellauitque do- 16:31 And the house of Israel called the name thereof

Manna: and it was like coriander seed, white; and the taste of it was like wafers made with honey.

mus Israhel nomen eius man quod erat quasi semen coriandri album gustusque eius quasi similae cum melle 16:32 dixit 16:32 And Moses said, This is the thing which the LORD

commandeth, Fill an omer of it to be kept for your generations; that they may see the bread wherewith I have fed you in the wilderness, when I brought you forth from the land of Egypt.

autem Moses iste est sermo quem praecepit Dominus imple gomor ex eo et custodiatur in futuras retro generationes ut nouerint panem quo alui uos in solitudine quando educti estis de terra Aegypti 16:33 dixitque Moses ad Aaron sume 16:33 And Moses said unto Aaron, Take a pot, and put an

omer full of manna therein, and lay it up before the LORD, to be kept for your generations.

uas unum et mitte ibi man quantum potest capere gomor et repone coram Domino ad seruandum in generationes uestras 16:34 sicut praecepit Dominus Mosi posuitque illud Aaron in 16:34 As the LORD commanded Moses, so Aaron laid it

up before the Testimony, to be kept.

tabernaculo reseruandum 16:35 filii autem Israhel comede- 16:35 And the children of Israel did eat manna forty years,

until they came to a land inhabited; they did eat manna, until they came unto the borders of the land of Canaan.

runt man quadraginta annis donec uenirent in terram habi-

Biblia Sacra Vulgata

tabilem hoc cibo aliti sunt usquequo tangerent fines terrae

16:36 Now an omer is the tenth part of an ephah. Chanaan 16:36 gomor autem decima pars est oephi

17:1 And all the congregation of the children of Israel 17:1 igitur profecta omnis multitudo filiorum Israhel de

journeyed from the wilderness of Sin, after their journeys, according to the commandment of the LORD, and pitched in Rephidim: and there was no water for the people to drink.

deserto Sin per mansiones suas iuxta sermonem Domini castrametata est in Raphidim ubi non erat aqua ad bibendum

17:2 Wherefore the people did chide with Moses, and said, populo 17:2 qui iurgatus contra Mosen ait da nobis aquam Give us water that we may drink. And Moses said unto them, Why chide ye with me? wherefore do ye tempt the

LORD? ut bibamus quibus respondit Moses quid iurgamini contra 17:3 And the people thirsted there for water; and the peo- me cur temptatis Dominum 17:3 sitiuit ergo populus ibi pro

ple murmured against Moses, and said, Wherefore is this that thou hast brought us up out of Egypt, to kill us and

our children and our cattle with thirst? aquae penuria et murmurauit contra Mosen dicens cur nos

exire fecisti de Aegypto ut occideres et nos et liberos no-

17:4 And Moses cried unto the LORD, saying, What shall stros ac iumenta siti 17:4 clamauit autem Moses ad Domi-

I do unto this people? they be almost ready to stone me.

num dicens quid faciam populo huic adhuc pauxillum et la-

17:5 And the LORD said unto Moses, Go on before the pidabunt me 17:5 ait Dominus ad Mosen antecede popu-

people, and take with thee of the elders of Israel; and thy rod, wherewith thou smotest the river, take in thine hand,

and go. lum et sume tecum de senibus Israhel et uirgam qua percus-

17:6 Behold, I will stand before thee there upon the rock in sisti fluuium tolle in manu tua et uade 17:6 en ego stabo Horeb; and thou shalt smite the rock, and there shall come water out of it, that the people may drink. And Moses did

so in the sight of the elders of Israel. coram te ibi super petram Horeb percutiesque petram et ex-

ibit ex ea aqua ut bibat populus fecit Moses ita coram se-

17:7 And he called the name of the place Massah, and nibus Israhel 17:7 et uocauit nomen loci illius Temptatio

Meribah, because of the chiding of the children of Israel, and because they tempted the LORD, saying, Is the LORD

among us, or not? propter iurgium filiorum Israhel et quia temptauerunt Do-17:8 Then came Amalek, and fought with Israel in minum dicentes estne Dominus in nobis an non 17:8 ue-

Rephidim.

nit autem Amalech et pugnabat contra Israhel in Raphidim

17:9 And Moses said unto Joshua, Choose us out men, and 17:9 dixitque Moses ad Iosue elige uiros et egressus pugna

go out, fight with Amalek: to morrow I will stand on the top of the hill with the rod of God in mine hand.

contra Amalech cras ego stabo in uertice collis habens uir-

17:10 So Joshua did as Moses had said to him, and fought gam Dei in manu mea 17:10 fecit Iosue ut locutus ei erat with Amalek: and Moses, Aaron, and Hur went up to the top of the hill.

Moses et pugnauit contra Amalech Moses autem et Aaron

17:11 And it came to pass, when Moses held up his hand, et Hur ascenderunt super uerticem collis 17:11 cumque le-

that Israel prevailed: and when he let down his hand, Amalek prevailed.

uaret Moses manus uincebat Israhel sin autem paululum re-

17:12 But Moses hands were heavy; and they took a stone, misisset superabat Amalech 17:12 manus autem Mosi erant

and put it under him, and he sat thereon; and Aaron and Hur stayed up his hands, the one on the one side, and the other on the other side; and his hands were steady until the going down of the sun.

graues sumentes igitur lapidem posuerunt subter eum in quo sedit Aaron autem et Hur sustentabant manus eius ex utraque parte et factum est ut manus ipsius non lassarentur usque ad

17:13 And Joshua discomfited Amalek and his people occasum solis 17:13 fugauitque Iosue Amalech et populum

with the edge of the sword.

17:14 And the LORD said unto Moses, Write this for a eius in ore gladii 17:14 dixit autem Dominus ad Mosen sc-

memorial in a book, and rehearse it in the ears of Joshua: for I will utterly put out the remembrance of Amalek from under heaven.

ribe hoc ob monumentum in libro et trade auribus Iosue delebo enim memoriam Amalech sub caelo 17:15 aedificauit- 17:15 And Moses built an altar, and called the name of it

Jehovahnissi:

que Moses altare et uocauit nomen eius Dominus exaltatio mea dicens 17:16 quia manus solii Domini et bellum Dei 17:16 For he said, Because the LORD hath sworn that the

LORD will have war with Amalek from generation to generation.

erit contra Amalech a generatione in generationem

18:1 cumque audisset Iethro sacerdos Madian cognatus 18:1 When Jethro, the priest of Midian, Moses’ father in law, heard of all that God had done for Moses, and for Israel his people, and that the LORD had brought Israel

Mosi omnia quae fecerat Deus Mosi et Israhel populo suo out of Egypt; eo quod eduxisset Dominus Israhel de Aegypto 18:2 tulit 18:2 Then Jethro, Moses’ father in law, took Zipporah,

Moses’ wife, after he had sent her back,

Sefforam uxorem Mosi quam remiserat 18:3 et duos filios 18:3 And her two sons; of which the name of the one was

Gershom; for he said, I have been an alien in a strange land:

eius quorum unus uocabatur Gersan dicente patre aduena fui in terra aliena 18:4 alter uero Eliezer Deus enim ait patris 18:4 And the name of the other was Eliezer; for the God of

my father, said he, was mine help, and delivered me from the sword of Pharaoh:

mei adiutor meus et eruit me de gladio Pharaonis 18:5 ue-

18:5 And Jethro, Moses’ father in law, came with his sons and his wife unto Moses into the wilderness, where he encamped at the mount of God:

nit ergo Iethro cognatus Mosi et filii eius et uxor ad Mosen in desertum ubi erat castrametatus iuxta montem Dei 18:6 18:6 And he said unto Moses, I thy father in law Jethro am

come unto thee, and thy wife, and her two sons with her.

et mandauit Mosi dicens ego cognatus tuus Iethro uenio ad te et uxor tua et duo filii tui cum ea 18:7 qui egressus in 18:7 And Moses went out to meet his father in law, and

did obeisance, and kissed him; and they asked each other of their welfare; and they came into the tent.

occursum cognati sui adorauit et osculatus est eum salutaueruntque se mutuo uerbis pacificis cumque intrasset tabernaculum 18:8 narrauit Moses cognato suo cuncta quae fecerat 18:8 And Moses told his father in law all that the LORD

had done unto Pharaoh and to the Egyptians for Israel’s sake, and all the travail that had come upon them by the

Deus Pharaoni et Aegyptiis propter Israhel uniuersum labo- way, and how the LORD delivered them. rem qui accidisset eis in itinere quo liberarat eos Dominus 18:9 laetatusque est Iethro super omnibus bonis quae fecerat 18:9 And Jethro rejoiced for all the goodness which the

LORD had done to Israel, whom he had delivered out of the hand of the Egyptians.

Dominus Israheli eo quod eruisset eum de manu Aegyptiorum 18:10 et ait benedictus Dominus qui liberauit uos de 18:10 And Jethro said, Blessed be the LORD, who hath

delivered you out of the hand of the Egyptians, and out of the hand of Pharaoh, who hath delivered the people from

manu Aegyptiorum et de manu Pharaonis qui eruit populum under the hand of the Egyptians.

suum de manu Aegypti 18:11 nunc cognoui quia magnus 18:11 Now I know that the LORD is greater than all gods: for in the thing wherein they dealt proudly he was above them.

Dominus super omnes deos eo quod superbe egerint contra illos 18:12 obtulit ergo Iethro cognatus Mosi holocausta 18:12 And Jethro, Moses’ father in law, took a burnt of-

fering and sacrifices for God: and Aaron came, and all the elders of Israel, to eat bread with Moses’ father in law

et hostias Deo ueneruntque Aaron et omnes senes Israhel before God. ut comederent panem cum eo coram Domino 18:13 altero 18:13 And it came to pass on the morrow, that Moses sat

to judge the people: and the people stood by Moses from the morning unto the evening.

autem die sedit Moses ut iudicaret populum qui adsistebat

Mosi de mane usque ad uesperam 18:14 quod cum uidis- 18:14 And when Moses’ father in law saw all that he did to the people, he said, What is this thing that thou doest to the people? why sittest thou thyself alone, and all the

set cognatus eius omnia scilicet quae agebat in populo ait people stand by thee from morning unto even?

Biblia Sacra Vulgata

quid est hoc quod facis in plebe cur solus sedes et omnis

18:15 And Moses said unto his father in law, Because the populus praestolatur de mane usque ad uesperam 18:15 cui

people come unto me to enquire of God:

respondit Moses uenit ad me populus quaerens sententiam

18:16 When they have a matter, they come unto me; and I Dei 18:16 cumque acciderit eis aliqua disceptatio ueniunt

judge between one and another, and I do make them know the statutes of God, and his laws.

ad me ut iudicem inter eos et ostendam praecepta Dei et

18:17 And Moses’ father in law said unto him, The thing leges eius 18:17 at ille non bonam inquit rem facis 18:18

that thou doest is not good. 18:18 Thou wilt surely wear away, both thou, and this people that is with thee: for this thing is too heavy for thee; thou art not able to perform it thyself alone.

stulto labore consumeris et tu et populus iste qui tecum est ultra uires tuas est negotium solus illud non poteris susti-

18:19 Hearken now unto my voice, I will give thee coun- nere 18:19 sed audi uerba mea atque consilia et erit Deus

sel, and God shall be with thee: Be thou for the people to God-ward, that thou mayest bring the causes unto God:

tecum esto tu populo in his quae ad Deum pertinent ut refe-

18:20 And thou shalt teach them ordinances and laws, and ras quae dicuntur ad eum 18:20 ostendasque populo caeri-

shalt shew them the way wherein they must walk, and the work that they must do.

monias et ritum colendi uiamque per quam ingredi debeant

18:21 Moreover thou shalt provide out of all the people et opus quod facere 18:21 prouide autem de omni plebe ui-

able men, such as fear God, men of truth, hating covetousness; and place such over them, to be rulers of thousands, and rulers of hundreds, rulers of fifties, and rulers of tens: ros potentes et timentes Deum in quibus sit ueritas et qui

oderint auaritiam et constitue ex eis tribunos et centuriones

18:22 And let them judge the people at all seasons: and et quinquagenarios et decanos 18:22 qui iudicent populum

it shall be, that every great matter they shall bring unto thee, but every small matter they shall judge: so shall it be easier for thyself, and they shall bear the burden with thee. omni tempore quicquid autem maius fuerit referant ad te et

ipsi minora tantummodo iudicent leuiusque tibi sit partito in

18:23 If thou shalt do this thing, and God command thee alios onere 18:23 si hoc feceris implebis imperium Dei et

so, then thou shalt be able to endure, and all this people shall also go to their place in peace.

praecepta eius poteris sustentare et omnis hic populus reuer-

18:24 So Moses hearkened to the voice of his father in law, tetur cum pace ad loca sua 18:24 quibus auditis Moses fecit

and did all that he had said.

18:25 And Moses chose able men out of all Israel, and omnia quae ille suggesserat 18:25 et electis uiris strenuis

made them heads over the people, rulers of thousands, rulers of hundreds, rulers of fifties, and rulers of tens.

de cuncto Israhel constituit eos principes populi tribunos et

18:26 And they judged the people at all seasons: the hard centuriones et quinquagenarios et decanos 18:26 qui iudi-

causes they brought unto Moses, but every small matter they judged themselves.

cabant plebem omni tempore quicquid autem grauius erat

18:27 And Moses let his father in law depart; and he went referebant ad eum faciliora tantummodo iudicantes 18:27

his way into his own land.

dimisitque cognatum qui reuersus abiit in terram suam

19:1 In the third month, when the children of Israel were 19:1 mense tertio egressionis Israhel de terra Aegypti in

gone forth out of the land of Egypt, the same day came they into the wilderness of Sinai.

die hac uenerunt in solitudinem Sinai 19:2 nam profecti de

19:2 For they were departed from Rephidim, and were come to the desert of Sinai, and had pitched in the wilderness; and there Israel camped before the mount.

Raphidim et peruenientes usque in desertum Sinai castrametati sunt in eodem loco ibique Israhel fixit tentoria e re-

19:3 And Moses went up unto God, and the LORD called gione montis 19:3 Moses autem ascendit ad Deum uocauit-

unto him out of the mountain, saying, Thus shalt thou say to the house of Jacob, and tell the children of Israel;

que eum Dominus de monte et ait haec dices domui Iacob et adnuntiabis filiis Israhel 19:4 uos ipsi uidistis quae fecerim 19:4 Ye have seen what I did unto the Egyptians, and how

I bare you on eagles’ wings, and brought you unto myself.

Aegyptiis quomodo portauerim uos super alas aquilarum et adsumpserim mihi 19:5 si ergo audieritis uocem meam et 19:5 Now therefore, if ye will obey my voice indeed, and

keep my covenant, then ye shall be a peculiar treasure unto me above all people: for all the earth is mine:

custodieritis pactum meum eritis mihi in peculium de cunctis populis mea est enim omnis terra 19:6 et uos eritis mihi 19:6 And ye shall be unto me a kingdom of priests, and an

holy nation. These are the words which thou shalt speak unto the children of Israel.

regnum sacerdotale et gens sancta haec sunt uerba quae loqueris ad filios Israhel 19:7 uenit Moses et conuocatis ma- 19:7 And Moses came and called for the elders of the peo-

ple, and laid before their faces all these words which the LORD commanded him.

ioribus natu populi exposuit omnes sermones quos mandauerat Dominus 19:8 responditque uniuersus populus simul 19:8 And all the people answered together, and said, All

that the LORD hath spoken we will do. And Moses returned the words of the people unto the LORD.

cuncta quae locutus est Dominus faciemus cumque rettulisset Moses uerba populi ad Dominum 19:9 ait ei Dominus 19:9 And the LORD said unto Moses, Lo, I come unto thee

in a thick cloud, that the people may hear when I speak with thee, and believe thee for ever. And Moses told the

iam nunc ueniam ad te in caligine nubis ut audiat me po- words of the people unto the LORD. pulus loquentem ad te et credat tibi in perpetuum nuntiauit ergo Moses uerba populi ad Dominum 19:10 qui dixit ei 19:10 And the LORD said unto Moses, Go unto the peo-

ple, and sanctify them to day and to morrow, and let them wash their clothes,

uade ad populum et sanctifica illos hodie et cras lauentque uestimenta sua 19:11 et sint parati in diem tertium die enim 19:11 And be ready against the third day: for the third day

the LORD will come down in the sight of all the people upon mount Sinai.

tertio descendet Dominus coram omni plebe super montem Sinai 19:12 constituesque terminos populo per circuitum et 19:12 And thou shalt set bounds unto the people round

about, saying, Take heed to yourselves, that ye go not up into the mount, or touch the border of it: whosoever

dices cauete ne ascendatis in montem nec tangatis fines illius toucheth the mount shall be surely put to death:

omnis qui tetigerit montem morte morietur 19:13 manus 19:13 There shall not an hand touch it, but he shall surely be stoned, or shot through; whether it be beast or man, it shall not live: when the trumpet soundeth long, they shall

non tanget eum sed lapidibus opprimetur aut confodietur ia- come up to the mount. culis siue iumentum fuerit siue homo non uiuet cum coeperit clangere bucina tunc ascendant in montem 19:14 descendit- 19:14 And Moses went down from the mount unto the

people, and sanctified the people; and they washed their clothes.

que Moses de monte ad populum et sanctificauit eum cumque lauissent uestimenta sua 19:15 ait ad eos estote parati in 19:15 And he said unto the people, Be ready against the

third day: come not at your wives.

diem tertium ne adpropinquetis uxoribus uestris 19:16 iam 19:16 And it came to pass on the third day in the morning,

that there were thunders and lightnings, and a thick cloud upon the mount, and the voice of the trumpet exceeding

aduenerat tertius dies et mane inclaruerat et ecce coeperunt loud; so that all the people that was in the camp trembled. audiri tonitrua ac micare fulgura et nubes densissima operire montem clangorque bucinae uehementius perstrepebat timuit populus qui erat in castris 19:17 cumque eduxisset 19:17 And Moses brought forth the people out of the camp

to meet with God; and they stood at the nether part of the mount.

eos Moses in occursum Dei de loco castrorum steterunt ad radices montis 19:18 totus autem mons Sinai fumabat eo 19:18 And mount Sinai was altogether on a smoke, be-

cause the LORD descended upon it in fire: and the smoke thereof ascended as the smoke of a furnace, and the whole

quod descendisset Dominus super eum in igne et ascenderet mount quaked greatly.

Biblia Sacra Vulgata

fumus ex eo quasi de fornace eratque mons omnis terribi-

19:19 And when the voice of the trumpet sounded long, lis 19:19 et sonitus bucinae paulatim crescebat in maius et

and waxed louder and louder, Moses spake, and God answered him by a voice.

prolixius tendebatur Moses loquebatur et Dominus respon-

19:20 And the LORD came down upon mount Sinai, on debat ei 19:20 descenditque Dominus super montem Sinai

the top of the mount: and the LORD called Moses up to the top of the mount; and Moses went up.

in ipso montis uertice et uocauit Mosen in cacumen eius quo

19:21 And the LORD said unto Moses, Go down, charge cum ascendisset 19:21 dixit ad eum descende et contestare

the people, lest they break through unto the LORD to gaze, and many of them perish.

populum ne forte uelint transcendere terminos ad uidendum

19:22 And let the priests also, which come near to the Dominum et pereat ex eis plurima multitudo 19:22 sacer-

LORD, sanctify themselves, lest the LORD break forth upon them.

dotes quoque qui accedunt ad Dominum sanctificentur ne

19:23 And Moses said unto the LORD, The people cannot percutiat eos 19:23 dixitque Moses ad Dominum non pote-

come up to mount Sinai: for thou chargedst us, saying, Set bounds about the mount, and sanctify it.

rit uulgus ascendere in montem Sinai tu enim testificatus es et iussisti dicens pone terminos circa montem et sanctifica il-

19:24 And the LORD said unto him, Away, get thee down, lum 19:24 cui ait Dominus uade descende ascendesque tu et

and thou shalt come up, thou, and Aaron with thee: but let not the priests and the people break through to come up

unto the LORD, lest he break forth upon them. Aaron tecum sacerdotes autem et populus ne transeant ter-

minos nec ascendant ad Dominum ne forte interficiat illos

19:25 So Moses went down unto the people, and spake 19:25 descendit Moses ad populum et omnia narrauit eis

unto them.

20:1 And God spake all these words, saying, 20:1 locutus quoque est Dominus cunctos sermones hos 20:2 I am the LORD thy God, which have brought thee 20:2 ego sum Dominus Deus tuus qui eduxi te de terra Ae-

out of the land of Egypt, out of the house of bondage.

20:3 Thou shalt have no other gods before me. gypti de domo seruitutis 20:3 non habebis deos alienos co-20:4 Thou shalt not make unto thee any graven image, or ram me 20:4 non facies tibi sculptile neque omnem simili-

any likeness of any thing that is in heaven above, or that is in the earth beneath, or that is in the water under the earth.

tudinem quae est in caelo desuper et quae in terra deorsum

20:5 Thou shalt not bow down thyself to them, nor serve nec eorum quae sunt in aquis sub terra 20:5 non adorabis ea

them: for I the LORD thy God am a jealous God, visiting the iniquity of the fathers upon the children unto the third

and fourth generation of them that hate me; neque coles ego sum Dominus Deus tuus fortis zelotes uisi-

tans iniquitatem patrum in filiis in tertiam et quartam gene-

20:6 And shewing mercy unto thousands of them that love rationem eorum qui oderunt me 20:6 et faciens misericor-

me, and keep my commandments.

diam in milia his qui diligunt me et custodiunt praecepta mea

20:7 Thou shalt not take the name of the LORD thy God in 20:7 non adsumes nomen Domini Dei tui in uanum nec enim

vain; for the LORD will not hold him guiltless that taketh his name in vain.

habebit insontem Dominus eum qui adsumpserit nomen Do-

20:8 Remember the sabbath day, to keep it holy. mini Dei sui frustra 20:8 memento ut diem sabbati sancti-20:9 Six days shalt thou labour, and do all thy work: fices 20:9 sex diebus operaberis et facies omnia opera tua 20:10 But the seventh day is the sabbath of the LORD thy 20:10 septimo autem die sabbati Domini Dei tui non facies

God: in it thou shalt not do any work, thou, nor thy son, nor thy daughter, thy manservant, nor thy maidservant, nor

thy cattle, nor thy stranger that is within thy gates: omne opus tu et filius tuus et filia tua seruus tuus et ancilla 20:11 For in six days the LORD made heaven and earth, tua iumentum tuum et aduena qui est intra portas tuas 20:11

the sea, and all that in them is, and rested the seventh day: wherefore the LORD blessed the sabbath day, and hallowed it.

sex enim diebus fecit Dominus caelum et terram et mare et omnia quae in eis sunt et requieuit in die septimo idcirco benedixit Dominus diei sabbati et sanctificauit eum 20:12 ho- 20:12 Honour thy father and thy mother: that thy days may

be long upon the land which the LORD thy God giveth thee.

nora patrem tuum et matrem tuam ut sis longeuus super terram quam Dominus Deus tuus dabit tibi 20:13 non occides 20:13 Thou shalt not kill. 20:14 non moechaberis 20:15 non furtum facies 20:16 non 20:14 Thou shalt not commit adultery.

20:15 Thou shalt not steal.

20:16 Thou shalt not bear false witness against thy neighbour.

loqueris contra proximum tuum falsum testimonium 20:17

20:17 Thou shalt not covet thy neighbour’s house, thou shalt not covet thy neighbour’s wife, nor his manservant, nor his maidservant, nor his ox, nor his ass, nor any thing that is thy neighbour’s.

non concupisces domum proximi tui nec desiderabis uxorem eius non seruum non ancillam non bouem non asinum nec omnia quae illius sunt 20:18 cunctus autem populus uidebat 20:18 And all the people saw the thunderings, and the

lightnings, and the noise of the trumpet, and the mountain smoking: and when the people saw it, they removed,

uoces et lampadas et sonitum bucinae montemque fumantem and stood afar off. et perterriti ac pauore concussi steterunt procul 20:19 di- 20:19 And they said unto Moses, Speak thou with us, and

we will hear: but let not God speak with us, lest we die.

centes Mosi loquere tu nobis et audiemus non loquatur nobis Dominus ne forte moriamur 20:20 et ait Moses ad populum 20:20 And Moses said unto the people, Fear not: for God

is come to prove you, and that his fear may be before your faces, that ye sin not.

nolite timere ut enim probaret uos uenit Deus et ut terror illius esset in uobis et non peccaretis 20:21 stetitque populus 20:21 And the people stood afar off, and Moses drew near

unto the thick darkness where God was.

de longe Moses autem accessit ad caliginem in qua erat Deus 20:22 dixit praeterea Dominus ad Mosen haec dices filiis 20:22 And the LORD said unto Moses, Thus thou shalt

say unto the children of Israel, Ye have seen that I have talked with you from heaven.

Israhel uos uidistis quod de caelo locutus sum uobis 20:23

20:23 Ye shall not make with me gods of silver, neither shall ye make unto you gods of gold.

non facietis mecum deos argenteos nec deos aureos facietis uobis 20:24 altare de terra facietis mihi et offeretis super eo 20:24 An altar of earth thou shalt make unto me, and shalt

sacrifice thereon thy burnt offerings, and thy peace offerings, thy sheep, and thine oxen: in all places where I record my name I will come unto thee, and I will bless thee.

holocausta et pacifica uestra oues uestras et boues in omni loco in quo memoria fuerit nominis mei ueniam ad te et benedicam tibi 20:25 quod si altare lapideum feceris mihi non 20:25 And if thou wilt make me an altar of stone, thou

shalt not build it of hewn stone: for if thou lift up thy tool upon it, thou hast polluted it.

aedificabis illud de sectis lapidibus si enim leuaueris cultrum tuum super eo polluetur 20:26 non ascendes per gradus ad 20:26 Neither shalt thou go up by steps unto mine altar,

that thy nakedness be not discovered thereon.

altare meum ne reueletur turpitudo tua

21:1 haec sunt iudicia quae propones eis 21:2 si emeris 21:1 Now these are the judgments which thou shalt set

before them.

21:2 If thou buy an Hebrew servant, six years he shall serve: and in the seventh he shall go out free for nothing.

seruum hebraeum sex annis seruiet tibi in septimo egredietur liber gratis 21:3 cum quali ueste intrauerit cum tali exeat 21:3 If he came in by himself, he shall go out by himself:

if he were married, then his wife shall go out with him.

si habens uxorem et uxor egredietur simul 21:4 sin autem 21:4 If his master have given him a wife, and she have born

him sons or daughters; the wife and her children shall be her master’s, and he shall go out by himself.

dominus dederit illi uxorem et peperit filios et filias mulier et liberi eius erunt domini sui ipse uero exibit cum uestitu

Biblia Sacra Vulgata

21:5 And if the servant shall plainly say, I love my master, suo 21:5 quod si dixerit seruus diligo dominum meum et my wife, and my children; I will not go out free:

21:6 Then his master shall bring him unto the judges; he uxorem ac liberos non egrediar liber 21:6 offeret eum do-

shall also bring him to the door, or unto the door post; and his master shall bore his ear through with an aul; and he

shall serve him for ever. minus diis et adplicabitur ad ostium et postes perforabitque 21:7 And if a man sell his daughter to be a maidservant, aurem eius subula et erit ei seruus in saeculum 21:7 si quis

she shall not go out as the menservants do.

uendiderit filiam suam in famulam non egredietur sicut an-

21:8 If she please not her master, who hath betrothed her cillae exire consuerunt 21:8 si displicuerit oculis domini sui

to himself, then shall he let her be redeemed: to sell her unto a strange nation he shall have no power, seeing he

hath dealt deceitfully with her. cui tradita fuerit dimittet eam populo autem alieno uendendi 21:9 And if he have betrothed her unto his son, he shall non habet potestatem si spreuerit eam 21:9 sin autem filio

deal with her after the manner of daughters.

21:10 If he take him another wife; her food, her raiment, suo desponderit eam iuxta morem filiarum faciet illi 21:10

and her duty of marriage, shall he not diminish.

quod si alteram ei acceperit prouidebit puellae nuptias et ue-

21:11 And if he do not these three unto her, then shall she stimenta et pretium pudicitiae non negabit 21:11 si tria ista

go out free without money.

21:12 He that smiteth a man, so that he die, shall be surely non fecerit egredietur gratis absque pecunia 21:12 qui per-

put to death.

21:13 And if a man lie not in wait, but God deliver him cusserit hominem uolens occidere morte moriatur 21:13 qui

into his hand; then I will appoint thee a place whither he shall flee.

autem non est insidiatus sed Deus illum tradidit in manu eius

21:14 But if a man come presumptuously upon his neigh- constituam tibi locum quo fugere debeat 21:14 si quis de

bour, to slay him with guile; thou shalt take him from mine altar, that he may die.

industria occiderit proximum suum et per insidias ab altari

21:15 And he that smiteth his father, or his mother, shall meo euelles eum ut moriatur 21:15 qui percusserit patrem be surely put to death.

21:16 And he that stealeth a man, and selleth him, or if he suum et matrem morte moriatur 21:16 qui furatus fuerit ho-

be found in his hand, he shall surely be put to death.

minem et uendiderit eum conuictus noxae morte moriatur

21:17 And he that curseth his father, or his mother, shall 21:17 qui maledixerit patri suo et matri morte moriatur 21:18

surely be put to death. 21:18 And if men strive together, and one smite another with a stone, or with his fist, and he die not, but keepeth his bed:

si rixati fuerint uiri et percusserit alter proximum suum lapide uel pugno et ille mortuus non fuerit sed iacuerit in lec-

21:19 If he rise again, and walk abroad upon his staff, tulo 21:19 si surrexerit et ambulauerit foris super baculum

then shall he that smote him be quit: only he shall pay for the loss of his time, and shall cause him to be thoroughly

healed. suum innocens erit qui percussit ita tamen ut operas eius et 21:20 And if a man smite his servant, or his maid, with a inpensas in medicos restituat 21:20 qui percusserit seruum

rod, and he die under his hand; he shall be surely punished.

suum uel ancillam uirga et mortui fuerint in manibus eius

21:21 Notwithstanding, if he continue a day or two, he criminis reus erit 21:21 sin autem uno die superuixerit uel

shall not be punished: for he is his money.

21:22 If men strive, and hurt a woman with child, so that duobus non subiacebit poenae quia pecunia illius est 21:22

her fruit depart from her, and yet no mischief follow: he shall be surely punished, according as the woman’s husband will lay upon him; and he shall pay as the judges determine.

si rixati fuerint uiri et percusserit quis mulierem praegnantem et abortiuum quidem fecerit sed ipsa uixerit subiacebit damno quantum expetierit maritus mulieris et arbitri iudica-

21:23 And if any mischief follow, then thou shalt give life rint 21:23 sin autem mors eius fuerit subsecuta reddet ani-

for life, 21:24 Eye for eye, tooth for tooth, hand for hand, foot for mam pro anima 21:24 oculum pro oculo dentem pro defoot, nte manum pro manu pedem pro pede 21:25 adustionem 21:25 Burning for burning, wound for wound, stripe for

stripe.

pro adustione uulnus pro uulnere liuorem pro liuore 21:26 21:26 And if a man smite the eye of his servant, or the eye

of his maid, that it perish; he shall let him go free for his eye’s sake.

si percusserit quispiam oculum serui sui aut ancillae et luscos eos fecerit dimittet liberos pro oculo quem eruit 21:27 21:27 And if he smite out his manservant’s tooth, or his

maidservant’s tooth; he shall let him go free for his tooth’s sake.

dentem quoque si excusserit seruo uel ancillae suae similiter dimittet eos liberos 21:28 si bos cornu petierit uirum 21:28 If an ox gore a man or a woman, that they die: then

the ox shall be surely stoned, and his flesh shall not be eaten; but the owner of the ox shall be quit.

aut mulierem et mortui fuerint lapidibus obruetur et non comedentur carnes eius dominusque bouis innocens erit 21:29 21:29 But if the ox were wont to push with his horn in

time past, and it hath been testified to his owner, and he hath not kept him in, but that he hath killed a man or a woman; the ox shall be stoned, and his owner also shall be put to death.

quod si bos cornipeta fuerit ab heri et nudius tertius et contestati sunt dominum eius nec reclusit eum occideritque uirum aut mulierem et bos lapidibus obruetur et dominum illius occident 21:30 quod si pretium ei fuerit inpositum dabit 21:30 If there be laid on him a sum of money, then he shall

give for the ransom of his life whatsoever is laid upon him.

pro anima sua quicquid fuerit postulatus 21:31 filium quo- 21:31 Whether he have gored a son, or have gored a

daughter, according to this judgment shall it be done unto him.

que et filiam si cornu percusserit simili sententiae subiacebit 21:32 si seruum ancillamque inuaserit triginta siclos argenti 21:32 If the ox shall push a manservant or a maidservant;

he shall give unto their master thirty shekels of silver, and the ox shall be stoned.

dabit domino bos uero lapidibus opprimetur 21:33 si quis

21:33 And if a man shall open a pit, or if a man shall dig a pit, and not cover it, and an ox or an ass fall therein;

aperuerit cisternam et foderit et non operuerit eam cecideritque bos uel asinus in eam 21:34 dominus cisternae red- 21:34 The owner of the pit shall make it good, and give

money unto the owner of them; and the dead beast shall be his.

det pretium iumentorum quod autem mortuum est ipsius erit 21:35 si bos alienus bouem alterius uulnerarit et ille mortuus 21:35 And if one man’s ox hurt another’s, that he die; then

they shall sell the live ox, and divide the money of it; and the dead ox also they shall divide.

fuerit uendent bouem uiuum et diuident pretium cadauer autem mortui inter se dispertient 21:36 sin autem sciebat quod 21:36 Or if it be known that the ox hath used to push in

time past, and his owner hath not kept him in; he shall surely pay ox for ox; and the dead shall be his own.

bos cornipeta esset ab heri et nudius tertius et non custodiuit eum dominus suus reddet bouem pro boue et cadauer integrum accipiet

22:1 si quis furatus fuerit bouem aut ouem et occiderit uel 22:1 If a man shall steal an ox, or a sheep, and kill it, or

sell it; he shall restore five oxen for an ox, and four sheep for a sheep.

uendiderit quinque boues pro uno boue restituet et quattuor oues pro una oue 22:2 si effringens fur domum siue suffo- 22:2 If a thief be found breaking up, and be smitten that

he die, there shall no blood be shed for him.

diens fuerit inuentus et accepto uulnere mortuus fuerit percussor non erit reus sanguinis 22:3 quod si orto sole hoc 22:3 If the sun be risen upon him, there shall be blood

shed for him; for he should make full restitution; if he have nothing, then he shall be sold for his theft.

fecerit homicidium perpetrauit et ipse morietur si non habuerit quod pro furto reddat uenundabitur 22:4 si inuentum 22:4 If the theft be certainly found in his hand alive,

whether it be ox, or ass, or sheep; he shall restore double.

fuerit apud eum quod furatus est uiuens siue bos siue asi-

Biblia Sacra Vulgata

22:5 If a man shall cause a field or vineyard to be eaten, nus siue ouis duplum restituet 22:5 si laeserit quispiam ag-

and shall put in his beast, and shall feed in another man’s field; of the best of his own field, and of the best of his

own vineyard, shall he make restitution. rum uel uineam et dimiserit iumentum suum ut depascatur

aliena quicquid optimum habuerit in agro suo uel in uinea

22:6 If fire break out, and catch in thorns, so that the stacks pro damni aestimatione restituet 22:6 si egressus ignis in-

of corn, or the standing corn, or the field, be consumed therewith; he that kindled the fire shall surely make resti-

tution. uenerit spinas et conprehenderit aceruos frugum siue stan-

tes segetes in agris reddet damnum qui ignem succenderit

22:7 If a man shall deliver unto his neighbour money or 22:7 si quis commendauerit amico pecuniam aut uas in custo-

stuff to keep, and it be stolen out of the man’s house; if the thief be found, let him pay double.

diam et ab eo qui susceperat furto ablata fuerint si inuenitur

22:8 If the thief be not found, then the master of the house fur duplum reddet 22:8 si latet dominus domus adplicabitur

shall be brought unto the judges, to see whether he have put his hand unto his neighbour’s goods.

ad deos et iurabit quod non extenderit manum in rem pro-

22:9 For all manner of trespass, whether it be for ox, for ximi sui 22:9 ad perpetrandam fraudem tam in boue quam ass, for sheep, for raiment, or for any manner of lost thing which another challengeth to be his, the cause of both parties shall come before the judges; and whom the judges shall condemn, he shall pay double unto his neighbour.

in asino et oue ac uestimento et quicquid damnum inferre potest ad deos utriusque causa perueniet et si illi iudicaue-

22:10 If a man deliver unto his neighbour an ass, or an ox, rint duplum restituet proximo suo 22:10 si quis commenda-

or a sheep, or any beast, to keep; and it die, or be hurt, or driven away, no man seeing it:

uerit proximo suo asinum bouem ouem et omne iumentum ad custodiam et mortuum fuerit aut debilitatum uel captum

22:11 Then shall an oath of the LORD be between them ab hostibus nullusque hoc uiderit 22:11 iusiurandum erit in

both, that he hath not put his hand unto his neighbour’s goods; and the owner of it shall accept thereof, and he

shall not make it good. medio quod non extenderit manum ad rem proximi sui su-

scipietque dominus iuramentum et ille reddere non cogetur

22:12 And if it be stolen from him, he shall make restitu- 22:12 quod si furto ablatum fuerit restituet damnum domino

tion unto the owner thereof. 22:13 If it be torn in pieces, then let him bring it for wit- 22:13 si comestum a bestia deferet ad eum quod occisum est ness, and he shall not make good that which was torn.

22:14 And if a man borrow ought of his neighbour, and it et non restituet 22:14 qui a proximo suo quicquam horum

be hurt, or die, the owner thereof being not with it, he shall surely make it good.

mutuo postularit et debilitatum aut mortuum fuerit domino

22:15 But if the owner thereof be with it, he shall not make non praesente reddere conpelletur 22:15 quod si inpraesen-

it good: if it be an hired thing, it came for his hire.

tiarum fuit dominus non restituet maxime si conductum ue-

22:16 And if a man entice a maid that is not betrothed, and nerat pro mercede operis sui 22:16 si seduxerit quis uirgi-

lie with her, he shall surely endow her to be his wife.

nem necdum desponsatam et dormierit cum ea dotabit eam

22:17 If her father utterly refuse to give her unto him, he et habebit uxorem 22:17 si pater uirginis dare noluerit re-

shall pay money according to the dowry of virgins.

ddet pecuniam iuxta modum dotis quam uirgines accipere

22:18 Thou shalt not suffer a witch to live. consuerunt 22:18 maleficos non patieris uiuere 22:19 qui

22:19 Whosoever lieth with a beast shall surely be put to death.

coierit cum iumento morte moriatur 22:20 qui immolat diis

22:20 He that sacrificeth unto any god, save unto the LORD only, he shall be utterly destroyed.

occidetur praeter Domino soli 22:21 aduenam non contris-

22:21 Thou shalt neither vex a stranger, nor oppress him: for ye were strangers in the land of Egypt.

tabis neque adfliges eum aduenae enim et ipsi fuistis in terra Aegypti 22:22 uiduae et pupillo non nocebitis 22:23 si lae- 22:22 Ye shall not afflict any widow, or fatherless child.

22:23 If thou afflict them in any wise, and they cry at all unto me, I will surely hear their cry;

seritis eos uociferabuntur ad me et ego audiam clamorem eorum 22:24 et indignabitur furor meus percutiamque uos gla- 22:24 And my wrath shall wax hot, and I will kill you

with the sword; and your wives shall be widows, and your children fatherless.

dio et erunt uxores uestrae uiduae et filii uestri pupilli 22:25

22:25 If thou lend money to any of my people that is poor by thee, thou shalt not be to him as an usurer, neither shalt thou lay upon him usury.

si pecuniam mutuam dederis populo meo pauperi qui habitat tecum non urgues eum quasi exactor nec usuris opprimes 22:26 si pignus a proximo tuo acceperis uestimentum ante so- 22:26 If thou at all take thy neighbour’s raiment to pledge,

thou shalt deliver it unto him by that the sun goeth down:

lis occasum redde ei 22:27 ipsum enim est solum quo operi- 22:27 For that is his covering only, it is his raiment for his skin: wherein shall he sleep? and it shall come to pass, when he crieth unto me, that I will hear; for I am gracious.

tur indumentum carnis eius nec habet aliud in quo dormiat si

clamauerit ad me exaudiam eum quia misericors sum 22:28 22:28 Thou shalt not revile the gods, nor curse the ruler of

thy people.

diis non detrahes et principi populi tui non maledices 22:29 22:29 Thou shalt not delay to offer the first of thy ripe

fruits, and of thy liquors: the firstborn of thy sons shalt thou give unto me.

decimas tuas et primitias non tardabis offerre primogenitum

filiorum tuorum dabis mihi 22:30 de bubus quoque et oui- 22:30 Likewise shalt thou do with thine oxen, and with thy sheep: seven days it shall be with his dam; on the eighth day thou shalt give it me.

bus similiter facies septem diebus sit cum matre sua die octauo reddes illum mihi 22:31 uiri sancti eritis mihi carnem 22:31 And ye shall be holy men unto me: neither shall ye

eat any flesh that is torn of beasts in the field; ye shall cast it to the dogs.

quae a bestiis fuerit praegustata non comedetis sed proicietis canibus

23:1 non suscipies uocem mendacii nec iunges manum 23:1 Thou shalt not raise a false report: put not thine hand

with the wicked to be an unrighteous witness.

tuam ut pro impio dicas falsum testimonium 23:2 non se- 23:2 Thou shalt not follow a multitude to do evil; neither shalt thou speak in a cause to decline after many to wrest judgment:

queris turbam ad faciendum malum nec in iudicio plurimorum adquiesces sententiae ut a uero deuies 23:3 pauperis 23:3 Neither shalt thou countenance a poor man in his

cause.

quoque non misereberis in negotio 23:4 si occurreris boui 23:4 If thou meet thine enemy’s ox or his ass going astray,

thou shalt surely bring it back to him again.

inimici tui aut asino erranti reduc ad eum 23:5 si uideris 23:5 If thou see the ass of him that hateth thee lying under

his burden, and wouldest forbear to help him, thou shalt surely help with him.

asinum odientis te iacere sub onere non pertransibis sed subleuabis cum eo 23:6 non declinabis in iudicio pauperis 23:6 Thou shalt not wrest the judgment of thy poor in his

cause.

23:7 mendacium fugies insontem et iustum non occides quia 23:7 Keep thee far from a false matter; and the inno-

cent and righteous slay thou not: for I will not justify the wicked.

auersor impium 23:8 nec accipias munera quae excaecant

23:8 And thou shalt take no gift: for the gift blindeth the wise, and perverteth the words of the righteous.

etiam prudentes et subuertunt uerba iustorum 23:9 pereg- 23:9 Also thou shalt not oppress a stranger: for ye know

the heart of a stranger, seeing ye were strangers in the land of Egypt.

rino molestus non eris scitis enim aduenarum animas quia et ipsi peregrini fuistis in terra Aegypti 23:10 sex annis se- 23:10 And six years thou shalt sow thy land, and shalt

gather in the fruits thereof:

minabis terram tuam et congregabis fruges eius 23:11 anno 23:11 But the seventh year thou shalt let it rest and lie still; that the poor of thy people may eat: and what they leave the beasts of the field shall eat. In like manner thou shalt

autem septimo dimittes eam et requiescere facies ut come- deal with thy vineyard, and with thy oliveyard. dant pauperes populi tui et quicquid reliqui fuerit edant bes-

Biblia Sacra Vulgata

23:12 Six days thou shalt do thy work, and on the seventh tiae agri ita facies in uinea et in oliueto tuo 23:12 sex diebus

day thou shalt rest: that thine ox and thine ass may rest, and the son of thy handmaid, and the stranger, may be

refreshed. operaberis septima die cessabis ut requiescat bos et asinus 23:13 And in all things that I have said unto you be cir- tuus et refrigeretur filius ancillae tuae et aduena 23:13 om-

cumspect: and make no mention of the name of other gods, neither let it be heard out of thy mouth.

nia quae dixi uobis custodite et per nomen externorum de-

23:14 Three times thou shalt keep a feast unto me in the orum non iurabitis neque audietur ex ore uestro 23:14 tri-

year.

23:15 Thou shalt keep the feast of unleavened bread: (thou bus uicibus per singulos annos mihi festa celebrabitis 23:15

shalt eat unleavened bread seven days, as I commanded thee, in the time appointed of the month Abib; for in it thou camest out from Egypt: and none shall appear before me empty:)

sollemnitatem azymorum custodies septem diebus comedes azyma sicut praecepi tibi tempore mensis nouorum quando egressus es de Aegypto non apparebis in conspectu meo ua-

23:16 And the feast of harvest, the firstfruits of thy cuus 23:16 et sollemnitatem messis primitiuorum operis tui

labours, which thou hast sown in the field: and the feast of ingathering, which is in the end of the year, when thou

hast gathered in thy labours out of the field. quaecumque serueris in agro sollemnitatem quoque in ex-

itu anni quando congregaueris omnes fruges tuas de agro

23:17 Three items in the year all thy males shall appear 23:17 ter in anno apparebit omne masculinum tuum coram before the LORD God.

23:18 Thou shalt not offer the blood of my sacrifice with Domino Deo 23:18 non immolabis super fermento sangui-

leavened bread; neither shall the fat of my sacrifice remain until the morning.

nem uictimae meae nec remanebit adeps sollemnitatis meae

23:19 The first of the firstfruits of thy land thou shalt bring usque mane 23:19 primitias frugum terrae tuae deferes in into the house of the LORD thy God. Thou shalt not seethe a kid in his mother’s milk.

domum Domini Dei tui nec coques hedum in lacte matris

23:20 Behold, I send an Angel before thee, to keep thee suae 23:20 ecce ego mittam angelum meum qui praecedat in the way, and to bring thee into the place which I have prepared.

te et custodiat in uia et introducat ad locum quem paraui

23:21 Beware of him, and obey his voice, provoke him 23:21 obserua eum et audi uocem eius nec contemnendum

not; for he will not pardon your transgressions: for my name is in him.

putes quia non dimittet cum peccaueritis et est nomen meum

23:22 But if thou shalt indeed obey his voice, and do all in illo 23:22 quod si audieris uocem eius et feceris omnia

that I speak; then I will be an enemy unto thine enemies, and an adversary unto thine adversaries.

quae loquor inimicus ero inimicis tuis et adfligam adfligen-

23:23 For mine Angel shall go before thee, and bring thee tes te 23:23 praecedetque te angelus meus et introducet te

in unto the Amorites, and the Hittites, and the Perizzites, and the Canaanites, the Hivites, and the Jebusites: and I

will cut them off. ad Amorreum et Hettheum et Ferezeum Chananeumque et 23:24 Thou shalt not bow down to their gods, nor serve Eueum et Iebuseum quos ego contribo 23:24 non adorabis

them, nor do after their works: but thou shalt utterly overthrow them, and quite break down their images.

deos eorum nec coles eos non facies opera eorum sed de-

23:25 And ye shall serve the LORD your God, and he shall strues eos et confringes statuas eorum 23:25 seruietisque

bless thy bread, and thy water; and I will take sickness away from the midst of thee.

Domino Deo uestro ut benedicam panibus tuis et aquis et

23:26 There shall nothing cast their young, nor be barren, auferam infirmitatem de medio tui 23:26 non erit infecunda in thy land: the number of thy days I will fulfil.

nec sterilis in terra tua numerum dierum tuorum implebo

23:27 I will send my fear before thee, and will destroy all 23:27 terrorem meum mittam in praecursum tuum et occidam the people to whom thou shalt come, and I will make all thine enemies turn their backs unto thee.

omnem populum ad quem ingredieris cunctorumque inimicorum tuorum coram te terga uertam 23:28 emittens crab- 23:28 And I will send hornets before thee, which shall

drive out the Hivite, the Canaanite, and the Hittite, from before thee.

rones prius qui fugabunt Eueum et Chananeum et Hettheum antequam introeas 23:29 non eiciam eos a facie tua anno 23:29 I will not drive them out from before thee in one

year; lest the land become desolate, and the beast of the field multiply against thee.

uno ne terra in solitudinem rediiatur et crescant contra te be-

stiae 23:30 paulatim expellam eos de conspectu tuo donec 23:30 By little and little I will drive them out from before

thee, until thou be increased, and inherit the land.

augearis et possideas terram 23:31 ponam autem terminos 23:31 And I will set thy bounds from the Red sea even

unto the sea of the Philistines, and from the desert unto the river: for I will deliver the inhabitants of the land into

tuos a mari Rubro usque ad mare Palestinorum et a deserto your hand; and thou shalt drive them out before thee. usque ad Fluuium tradam manibus uestris habitatores terrae et eiciam eos de conspectu uestro 23:32 non inibis cum eis 23:32 Thou shalt make no covenant with them, nor with

their gods.

foedus nec cum diis eorum 23:33 non habitent in terra tua 23:33 They shall not dwell in thy land, lest they make thee sin against me: for if thou serve their gods, it will surely be a snare unto thee.

ne forte peccare te faciant in me si seruieris diis eorum quod tibi certo erit in scandalum

24:1 Mosi quoque dixit ascende ad Dominum tu et Aa- 24:1 And he said unto Moses, Come up unto the LORD,

thou, and Aaron, Nadab, and Abihu, and seventy of the elders of Israel; and worship ye afar off.

ron Nadab et Abiu et septuaginta senes ex Israhel et adorabitis procul 24:2 solusque Moses ascendet ad Dominum et 24:2 And Moses alone shall come near the LORD: but

they shall not come nigh; neither shall the people go up with him.

illi non adpropinquabunt nec populus ascendet cum eo 24:3

24:3 And Moses came and told the people all the words of the LORD, and all the judgments: and all the people answered with one voice, and said, All the words which the LORD hath said will we do.

uenit ergo Moses et narrauit plebi omnia uerba Domini atque iudicia responditque cunctus populus una uoce omnia uerba Domini quae locutus est faciemus 24:4 scripsit autem Mo- 24:4 And Moses wrote all the words of the LORD, and

rose up early in the morning, and builded an altar under the hill, and twelve pillars, according to the twelve tribes

ses uniuersos sermones Domini et mane consurgens aedifi- of Israel. cauit altare ad radices montis et duodecim titulos per duodecim tribus Israhel 24:5 misitque iuuenes de filiis Israhel 24:5 And he sent young men of the children of Israel,

which offered burnt offerings, and sacrificed peace offerings of oxen unto the LORD.

et obtulerunt holocausta immolaueruntque uictimas pacificas Domino uitulos 24:6 tulit itaque Moses dimidiam par- 24:6 And Moses took half of the blood, and put it in ba-

sons; and half of the blood he sprinkled on the altar.

tem sanguinis et misit in crateras partem autem residuam fudit super altare 24:7 adsumensque uolumen foederis legit 24:7 And he took the book of the covenant, and read in the

audience of the people: and they said, All that the LORD hath said will we do, and be obedient.

audiente populo qui dixerunt omnia quae locutus est Dominus faciemus et erimus oboedientes 24:8 ille uero sump- 24:8 And Moses took the blood, and sprinkled it on the

people, and said, Behold the blood of the covenant, which the LORD hath made with you concerning all these words.

tum sanguinem respersit in populum et ait hic est sanguis foederis quod pepigit Dominus uobiscum super cunctis sermonibus his 24:9 ascenderuntque Moses et Aaron Nadab et 24:9 Then went up Moses, and Aaron, Nadab, and Abihu,

and seventy of the elders of Israel:

Abiu et septuaginta de senioribus Israhel 24:10 et uiderunt 24:10 And they saw the God of Israel: and there was under

his feet as it were a paved work of a sapphire stone, and as it were the body of heaven in his clearness.

Deum Israhel sub pedibus eius quasi opus lapidis sapphi-

Biblia Sacra Vulgata

24:11 And upon the nobles of the children of Israel he laid rini et quasi caelum cum serenum est 24:11 nec super eos not his hand: also they saw God, and did eat and drink.

qui procul recesserant de filiis Israhel misit manum suam ui-

24:12 And the LORD said unto Moses, Come up to me deruntque Deum et comederunt ac biberunt 24:12 dixit au-

into the mount, and be there: and I will give thee tables of stone, and a law, and commandments which I have written;

that thou mayest teach them. tem Dominus ad Mosen ascende ad me in montem et esto

ibi daboque tibi tabulas lapideas et legem ac mandata quae

24:13 And Moses rose up, and his minister Joshua: and scripsi ut doceas eos 24:13 surrexerunt Moses et Iosue mi-

Moses went up into the mount of God.

24:14 And he said unto the elders, Tarry ye here for us, nister eius ascendensque Moses in montem Dei 24:14 se-

until we come again unto you: and, behold, Aaron and Hur are with you: if any man have any matters to do, let

him come unto them. nioribus ait expectate hic donec reuertamur ad uos habetis

Aaron et Hur uobiscum si quid natum fuerit quaestionis refe-

24:15 And Moses went up into the mount, and a cloud retis ad eos 24:15 cumque ascendisset Moses operuit nubes

covered the mount.

24:16 And the glory of the LORD abode upon mount montem 24:16 et habitauit gloria Domini super Sinai tegens

Sinai, and the cloud covered it six days: and the seventh day he called unto Moses out of the midst of the cloud.

illum nube sex diebus septimo autem die uocauit eum de me-

24:17 And the sight of the glory of the LORD was like dio caliginis 24:17 erat autem species gloriae Domini quasi

devouring fire on the top of the mount in the eyes of the children of Israel.

ignis ardens super uerticem montis in conspectu filiorum Is-

24:18 And Moses went into the midst of the cloud, and gat rahel 24:18 ingressusque Moses medium nebulae ascendit

him up into the mount: and Moses was in the mount forty days and forty nights.

in montem et fuit ibi quadraginta diebus et quadraginta noctibus

25:1 And the LORD spake unto Moses, saying, 25:1 locutusque est Dominus ad Mosen dicens 25:2 lo-

25:2 Speak unto the children of Israel, that they bring me an offering: of every man that giveth it willingly with his heart ye shall take my offering.

quere filiis Israhel ut tollant mihi primitias ab omni homine

25:3 And this is the offering which ye shall take of them; qui offert ultroneus accipietis eas 25:3 haec sunt autem

gold, and silver, and brass, 25:4 And blue, and purple, and scarlet, and fine linen, and quae accipere debetis aurum et argentum et aes 25:4 hyagoats’ hair,

cinthum et purpuram coccumque bis tinctum et byssum pi-

25:5 And rams’ skins dyed red, and badgers’ skins, and los caprarum 25:5 et pelles arietum rubricatas pelles ianthi-

shittim wood,

25:6 Oil for the light, spices for anointing oil, and for nas et ligna setthim 25:6 oleum ad luminaria concinnanda

sweet incense,

25:7 Onyx stones, and stones to be set in the ephod, and aromata in unguentum et thymiama boni odoris 25:7 lapi-

in the breastplate.

des onychinos et gemmas ad ornandum ephod ac rationale

25:8 And let them make me a sanctuary; that I may dwell 25:8 facientque mihi sanctuarium et habitabo in medio eorum

among them.

25:9 According to all that I shew thee, after the pattern 25:9 iuxta omnem similitudinem tabernaculi quod ostendam

of the tabernacle, and the pattern of all the instruments thereof, even so shall ye make it.

tibi et omnium uasorum in cultum eius sicque facietis illud

25:10 And they shall make an ark of shittim wood: two 25:10 arcam de lignis setthim conpingite cuius longitudo ha-

cubits and a half shall be the length thereof, and a cubit and a half the breadth thereof, and a cubit and a half the

height thereof. beat duos semis cubitos latitudo cubitum et dimidium alti-25:11 And thou shalt overlay it with pure gold, within and tudo cubitum similiter ac semissem 25:11 et deaurabis eam

without shalt thou overlay it, and shalt make upon it a crown of gold round about.

auro mundissimo intus et foris faciesque supra coronam auream per circuitum 25:12 et quattuor circulos aureos quos 25:12 And thou shalt cast four rings of gold for it, and put

them in the four corners thereof; and two rings shall be in the one side of it, and two rings in the other side of it.

pones per quattuor arcae angulos duo circuli sint in latere uno et duo in altero 25:13 facies quoque uectes de lignis 25:13 And thou shalt make staves of shittim wood, and

overlay them with gold.

setthim et operies eos auro 25:14 inducesque per circulos 25:14 And thou shalt put the staves into the rings by the

sides of the ark, that the ark may be borne with them.

qui sunt in arcae lateribus ut portetur in eis 25:15 qui sem- 25:15 The staves shall be in the rings of the ark: they shall

not be taken from it.

per erunt in circulis nec umquam extrahentur ab eis 25:16 25:16 And thou shalt put into the ark the testimony which

I shall give thee.

ponesque in arcam testificationem quam dabo tibi 25:17 fa- 25:17 And thou shalt make a mercy seat of pure gold: two

cubits and a half shall be the length thereof, and a cubit and a half the breadth thereof.

cies et propitiatorium de auro mundissimo duos cubitos et dimidium tenebit longitudo eius cubitum ac semissem latitudo 25:18 duos quoque cherubin aureos et productiles fa- 25:18 And thou shalt make two cherubims of gold, of

beaten work shalt thou make them, in the two ends of the mercy seat.

cies ex utraque parte oraculi 25:19 cherub unus sit in latere

25:19 And make one cherub on the one end, and the other cherub on the other end: even of the mercy seat shall ye make the cherubims on the two ends thereof.

uno et alter in altero 25:20 utrumque latus propitiatorii te-

25:20 And the cherubims shall stretch forth their wings on high, covering the mercy seat with their wings, and their faces shall look one to another; toward the mercy seat shall the faces of the cherubims be.

gant expandentes alas et operientes oraculum respiciantque se mutuo uersis uultibus in propitiatorium quo operienda est arca 25:21 in qua pones testimonium quod dabo tibi 25:22 25:21 And thou shalt put the mercy seat above upon the

ark; and in the ark thou shalt put the testimony that I shall give thee. 25:22 And there I will meet with thee, and I will commune with thee from above the mercy seat, from between the two cherubims which are upon the ark of the testimony, of all things which I will give thee in commandment unto the children of Israel.

inde praecipiam et loquar ad te supra propitiatorio scilicet ac medio duorum cherubin qui erunt super arcam testimonii cuncta quae mandabo per te filiis Israhel 25:23 facies 25:23 Thou shalt also make a table of shittim wood: two

cubits shall be the length thereof, and a cubit the breadth thereof, and a cubit and a half the height thereof.

et mensam de lignis setthim habentem duos cubitos longitudinis et in latitudine cubitum et in altitudine cubitum ac semissem 25:24 et inaurabis eam auro purissimo faciesque 25:24 And thou shalt overlay it with pure gold, and make

thereto a crown of gold round about.

illi labium aureum per circuitum 25:25 et ipsi labio coro- 25:25 And thou shalt make unto it a border of an hand

breadth round about, and thou shalt make a golden crown to the border thereof round about.

nam interrasilem altam quattuor digitis et super illam alteram coronam aureolam 25:26 quattuor quoque circulos au- 25:26 And thou shalt make for it four rings of gold, and

put the rings in the four corners that are on the four feet thereof.

reos praeparabis et pones eos in quattuor angulis eiusdem mensae per singulos pedes 25:27 subter coronam erunt cir- 25:27 Over against the border shall the rings be for places

of the staves to bear the table.

culi aurei ut mittantur uectes per eos et possit mensa portari 25:28 ipsosque uectes facies de lignis setthim et circumdabis 25:28 And thou shalt make the staves of shittim wood, and

overlay them with gold, that the table may be borne with them.

auro ad subuehendam mensam 25:29 parabis et acetabula

25:29 And thou shalt make the dishes thereof, and spoons thereof, and covers thereof, and bowls thereof, to cover withal: of pure gold shalt thou make them.

ac fialas turibula et cyatos in quibus offerenda sunt libamina ex auro purissimo 25:30 et pones super mensam panes pro- 25:30 And thou shalt set upon the table shewbread before

me alway.

positionis in conspectu meo semper 25:31 facies et cande- 25:31 And thou shalt make a candlestick of pure gold: of

beaten work shall the candlestick be made: his shaft, and his branches, his bowls, his knops, and his flowers, shall

labrum ductile de auro mundissimo hastile eius et calamos be of the same.

Biblia Sacra Vulgata

25:32 And six branches shall come out of the sides of it; scyphos et spherulas ac lilia ex ipso procedentia 25:32 sex

three branches of the candlestick out of the one side, and three branches of the candlestick out of the other side:

calami egredientur de lateribus tres ex uno latere et tres ex

25:33 Three bowls made like unto almonds, with a knop altero 25:33 tres scyphi quasi in nucis modum per calamos

and a flower in one branch; and three bowls made like almonds in the other branch, with a knop and a flower: so

in the six branches that come out of the candlestick. singulos spherulaque simul et lilium et tres similiter scyphi

instar nucis in calamo altero spherulaque et lilium hoc erit

25:34 And in the candlesticks shall be four bowls made opus sex calamorum qui producendi sunt de hastili 25:34

like unto almonds, with their knops and their flowers.

in ipso autem candelabro erunt quattuor scyphi in nucis mo-

25:35 And there shall be a knop under two branches of the dum spherulaeque per singulos et lilia 25:35 spherula sub

same, and a knop under two branches of the same, and a knop under two branches of the same, according to the six

branches that proceed out of the candlestick. duobus calamis per tria loca qui simul sex fiunt procedentes 25:36 Their knops and their branches shall be of the same: de hastili uno 25:36 et spherae igitur et calami ex ipso erunt

all it shall be one beaten work of pure gold.

25:37 And thou shalt make the seven lamps thereof: and uniuersa ductilia de auro purissimo 25:37 facies et lucernas

they shall light the lamps thereof, that they may give light over against it.

septem et pones eas super candelabrum ut luceant ex ad-

25:38 And the tongs thereof, and the snuffdishes thereof, uerso 25:38 emunctoria quoque et ubi quae emuncta sunt shall be of pure gold.

25:39 Of a talent of pure gold shall he make it, with all extinguantur fient de auro purissimo 25:39 omne pondus

these vessels.

candelabri cum uniuersis uasis suis habebit talentum auri

25:40 And look that thou make them after their pattern, mundissimi 25:40 inspice et fac secundum exemplar quod

which was shewed thee in the mount.

tibi in monte monstratum est

26:1 Moreover thou shalt make the tabernacle with ten 26:1 tabernaculum uero ita fiet decem cortinas de bysso

curtains of fine twined linen, and blue, and purple, and scarlet: with cherubims of cunning work shalt thou make

them. retorta et hyacintho ac purpura coccoque bis tincto uariatas 26:2 The length of one curtain shall be eight and twenty opere plumario facies 26:2 longitudo cortinae unius habe-

cubits, and the breadth of one curtain four cubits: and every one of the curtains shall have one measure.

bit uiginti octo cubitos latitudo quattuor cubitorum erit unius

26:3 The five curtains shall be coupled together one to an- mensurae fient uniuersa tentoria 26:3 quinque cortinae sibi

other; and other five curtains shall be coupled one to another.

iungentur mutuo et aliae quinque nexu simili coherebunt

26:4 And thou shalt make loops of blue upon the edge 26:4 ansulas hyacinthinas in lateribus ac summitatibus facies

of the one curtain from the selvedge in the coupling; and likewise shalt thou make in the uttermost edge of another

curtain, in the coupling of the second. cortinarum ut possint inuicem copulari 26:5 quinquagenas

26:5 Fifty loops shalt thou make in the one curtain, and fifty loops shalt thou make in the edge of the curtain that is in the coupling of the second; that the loops may take hold one of another.

ansulas cortina habebit in utraque parte ita insertas ut ansa

26:6 And thou shalt make fifty taches of gold, and couple contra ansam ueniat et altera alteri possit aptari 26:6 facies

the curtains together with the taches: and it shall be one tabernacle.

et quinquaginta circulos aureos quibus cortinarum uela iun-

26:7 And thou shalt make curtains of goats’ hair to be a genda sunt ut unum tabernaculum fiat 26:7 facies et saga ci-

covering upon the tabernacle: eleven curtains shalt thou make.

licina undecim ad operiendum tectum tabernaculi 26:8 lon-

26:8 The length of one curtain shall be thirty cubits, and the breadth of one curtain four cubits: and the eleven curtains shall be all of one measure.

gitudo sagi unius habebit triginta cubitos et latitudo quattuor

26:9 And thou shalt couple five curtains by themselves, aequa erit mensura sagorum omnium 26:9 e quibus quinque

and six curtains by themselves, and shalt double the sixth curtain in the forefront of the tabernacle.

iunges seorsum et sex sibi mutuo copulabis ita ut sextum sagum in fronte tecti duplices 26:10 facies et quinquaginta an- 26:10 And thou shalt make fifty loops on the edge of the

one curtain that is outmost in the coupling, and fifty loops in the edge of the curtain which coupleth the second.

sas in ora sagi unius ut coniungi cum altero queat et quinquaginta ansas in ora sagi alterius ut cum altero copuletur 26:11 26:11 And thou shalt make fifty taches of brass, and put

the taches into the loops, and couple the tent together, that it may be one.

quinquaginta fibulas aeneas quibus iungantur ansae et unum ex omnibus operimentum fiat 26:12 quod autem superfuerit 26:12 And the remnant that remaineth of the curtains of

the tent, the half curtain that remaineth, shall hang over the backside of the tabernacle.

in sagis quae parantur tecto id est unum sagum quod amplius est ex medietate eius operies posteriora tabernaculi 26:13 et 26:13 And a cubit on the one side, and a cubit on the other

side of that which remaineth in the length of the curtains of the tent, it shall hang over the sides of the tabernacle on

cubitus ex una parte pendebit et alter ex altera qui plus est in this side and on that side, to cover it. sagorum longitudine utrumque latus tabernaculi protegens 26:14 facies et operimentum aliud tecto de pellibus arietum 26:14 And thou shalt make a covering for the tent of rams’

skins dyed red, and a covering above of badgers’ skins.

rubricatis et super hoc rursum aliud operimentum de ianthinis pellibus 26:15 facies et tabulas stantes tabernaculi de 26:15 And thou shalt make boards for the tabernacle of

shittim wood standing up.

lignis setthim 26:16 quae singulae denos cubitos in longi- 26:16 Ten cubits shall be the length of a board, and a cubit

and a half shall be the breadth of one board.

tudine habeant et in latitudine singulos ac semissem 26:17 26:17 Two tenons shall there be in one board, set in order

one against another: thus shalt thou make for all the boards of the tabernacle.

in lateribus tabulae duae incastraturae fient quibus tabula alteri tabulae conectatur atque in hunc modum cunctae tabulae parabuntur 26:18 quarum uiginti erunt in latere meridiano 26:18 And thou shalt make the boards for the tabernacle,

twenty boards on the south side southward.

quod uergit ad austrum 26:19 quibus quadraginta bases ar- 26:19 And thou shalt make forty sockets of silver under

the twenty boards; two sockets under one board for his two tenons, and two sockets under another board for his

genteas fundes ut binae bases singulis tabulis per duos an- two tenons. gulos subiciantur 26:20 in latere quoque secundo taberna- 26:20 And for the second side of the tabernacle on the

north side there shall be twenty boards:

culi quod uergit ad aquilonem uiginti tabulae erunt 26:21 26:21 And their forty sockets of silver; two sockets under

one board, and two sockets under another board.

quadraginta habentes bases argenteas binae bases singulis tabulis subponentur 26:22 ad occidentalem uero plagam ta- 26:22 And for the sides of the tabernacle westward thou

shalt make six boards.

bernaculi facies sex tabulas 26:23 et rursum alias duas quae 26:23 And two boards shalt thou make for the corners of

the tabernacle in the two sides.

in angulis erigantur post tergum tabernaculi 26:24 eruntque 26:24 And they shall be coupled together beneath, and

they shall be coupled together above the head of it unto one ring: thus shall it be for them both; they shall be for

coniunctae a deorsum usque sursum et una omnes conpago the two corners. retinebit duabus quoque tabulis quae in angulis ponendae sunt similis iunctura seruabitur 26:25 et erunt simul tabu- 26:25 And they shall be eight boards, and their sockets of

silver, sixteen sockets; two sockets under one board, and two sockets under another board.

lae octo bases earum argenteae sedecim duabus basibus per unam tabulam supputatis 26:26 facies et uectes de lignis 26:26 And thou shalt make bars of shittim wood; five for

the boards of the one side of the tabernacle,

setthim quinque ad continendas tabulas in uno latere tabernaculi 26:27 et quinque alios in altero et eiusdem numeri ad 26:27 And five bars for the boards of the other side of the

tabernacle, and five bars for the boards of the side of the tabernacle, for the two sides westward.

occidentalem plagam 26:28 qui mittentur per medias tabu-

26:28 And the middle bar in the midst of the boards shall reach from end to end.

Biblia Sacra Vulgata

26:29 And thou shalt overlay the boards with gold, and las a summo usque ad summum 26:29 ipsasque tabulas demake their rings of gold for places for the bars: and thou shalt overlay the bars with gold.

aurabis et fundes eis anulos aureos per quos uectes tabulata

26:30 And thou shalt rear up the tabernacle according to contineant quos operies lamminis aureis 26:30 et eriges ta-

the fashion thereof which was shewed thee in the mount.

bernaculum iuxta exemplum quod tibi in monte monstratum

26:31 And thou shalt make a vail of blue, and purple, and est 26:31 facies et uelum de hyacintho et purpura coccoque

scarlet, and fine twined linen of cunning work: with cherubims shall it be made:

bis tincto et bysso retorta opere plumario et pulchra uarietate

26:32 And thou shalt hang it upon four pillars of shittim contextum 26:32 quod adpendes ante quattuor columnas de

wood overlaid with gold: their hooks shall be of gold, upon the four sockets of silver.

lignis setthim quae ipsae quidem deauratae erunt et habe-

26:33 And thou shalt hang up the vail under the taches, bunt capita aurea sed bases argenteas 26:33 inseretur autem

that thou mayest bring in thither within the vail the ark of the testimony: and the vail shall divide unto you between

the holy place and the most holy. uelum per circulos intra quod pones arcam testimonii et quo 26:34 And thou shalt put the mercy seat upon the ark of sanctuarium et sanctuarii sanctuaria diuidentur 26:34 pones

the testimony in the most holy place.

et propitiatorium super arcam testimonii in sancta sancto-

26:35 And thou shalt set the table without the vail, and rum 26:35 mensamque extra uelum et contra mensam canthe candlestick over against the table on the side of the tabernacle toward the south: and thou shalt put the table

on the north side. delabrum in latere tabernaculi meridiano mensa enim stabit 26:36 And thou shalt make an hanging for the door of the in parte aquilonis 26:36 facies et tentorium in introitu taber-

tent, of blue, and purple, and scarlet, and fine twined linen, wrought with needlework.

naculi de hyacintho et purpura coccoque bis tincto et bysso

26:37 And thou shalt make for the hanging five pillars of retorta opere plumarii 26:37 et quinque columnas deaurabis

shittim wood, and overlay them with gold, and their hooks shall be of gold: and thou shalt cast five sockets of brass

for them. lignorum setthim ante quas ducetur tentorium quarum erunt capita aurea et bases aeneae

27:1 And thou shalt make an altar of shittim wood, five 27:1 facies et altare de lignis setthim quod habebit quinque

cubits long, and five cubits broad; the altar shall be foursquare: and the height thereof shall be three cubits.

cubitos in longitudine et totidem in latitudine id est quad-

27:2 And thou shalt make the horns of it upon the four rum et tres cubitos in altitudine 27:2 cornua autem per quat-

corners thereof: his horns shall be of the same: and thou shalt overlay it with brass.

tuor angulos ex ipso erunt et operies illud aere 27:3 facies-

27:3 And thou shalt make his pans to receive his ashes, and his shovels, and his basons, and his fleshhooks, and his firepans: all the vessels thereof thou shalt make of brass.

que in usus eius lebetas ad suscipiendos cineres et forcipes atque fuscinulas et ignium receptacula omnia uasa ex aere

27:4 And thou shalt make for it a grate of network of brass; fabricabis 27:4 craticulamque in modum retis aeneam per and upon the net shalt thou make four brasen rings in the four corners thereof.

cuius quattuor angulos erunt quattuor anuli aenei 27:5 quos

27:5 And thou shalt put it under the compass of the altar beneath, that the net may be even to the midst of the altar.

pones subter arulam altaris eritque craticula usque ad alta-

27:6 And thou shalt make staves for the altar, staves of ris medium 27:6 facies et uectes altaris de lignis setthim

shittim wood, and overlay them with brass.

27:7 And the staves shall be put into the rings, and the duos quos operies lamminis aeneis 27:7 et induces per cir-

staves shall be upon the two sides of the altar, to bear it.

27:8 Hollow with boards shalt thou make it: as it was culos eruntque ex utroque latere altaris ad portandum 27:8

shewed thee in the mount, so shall they make it.

non solidum sed inane et cauum intrinsecus facies illud sicut

27:9 And thou shalt make the court of the tabernacle: for tibi in monte monstratum est 27:9 facies et atrium taberna-

the south side southward there shall be hangings for the court of fine twined linen of an hundred cubits long for one side:

culi in cuius plaga australi contra meridiem erunt tentoria de bysso retorta centum cubitos unum latus tenebit in longitudine 27:10 et columnas uiginti cum basibus totidem aeneis 27:10 And the twenty pillars thereof and their twenty

sockets shall be of brass; the hooks of the pillars and their fillets shall be of silver.

quae capita cum celaturis suis habebunt argentea 27:11 si-

27:11 And likewise for the north side in length there shall be hangings of an hundred cubits long, and his twenty pillars and their twenty sockets of brass; the hooks of the pillars and their fillets of silver.

militer in latere aquilonis per longum erunt tentoria centum cubitorum columnae uiginti et bases aeneae eiusdem numeri et capita earum cum celaturis suis argentea 27:12 in lati- 27:12 And for the breadth of the court on the west side

shall be hangings of fifty cubits: their pillars ten, and their sockets ten.

tudine uero atrii quod respicit ad occidentem erunt tentoria per quinquaginta cubitos et columnae decem basesque totidem 27:13 in ea quoque atrii latitudine quae respicit ad 27:13 And the breadth of the court on the east side east-

ward shall be fifty cubits.

orientem quinquaginta cubiti erunt 27:14 in quibus quinde- 27:14 The hangings of one side of the gate shall be fifteen

cubits: their pillars three, and their sockets three.

cim cubitorum tentoria lateri uno deputabuntur columnaeque tres et bases totidem 27:15 et in latere altero erunt tento- 27:15 And on the other side shall be hangings fifteen cu-

bits: their pillars three, and their sockets three.

ria cubitos obtinentia quindecim columnas tres et bases totidem 27:16 in introitu uero atrii fiet tentorium cubitorum 27:16 And for the gate of the court shall be an hanging

of twenty cubits, of blue, and purple, and scarlet, and fine twined linen, wrought with needlework: and their pillars

uiginti ex hyacintho et purpura coccoque bis tincto et bysso shall be four, and their sockets four. retorta opere plumarii columnas habebit quattuor cum basi-

bus totidem 27:17 omnes columnae atrii per circuitum ue- 27:17 All the pillars round about the court shall be filleted with silver; their hooks shall be of silver, and their sockets of brass.

stitae erunt argenti lamminis capitibus argenteis et basibus aeneis 27:18 in longitudine occupabit atrium cubitos cen- 27:18 The length of the court shall be an hundred cubits,

and the breadth fifty every where, and the height five cubits of fine twined linen, and their sockets of brass.

tum in latitudine quinquaginta altitudo quinque cubitorum erit fietque de bysso retorta et habebit bases aeneas 27:19 27:19 All the vessels of the tabernacle in all the service

thereof, and all the pins thereof, and all the pins of the court, shall be of brass.

cuncta uasa tabernaculi in omnes usus et caerimonias tam paxillos eius quam atrii ex aere facies 27:20 praecipe filiis 27:20 And thou shalt command the children of Israel, that

they bring thee pure oil olive beaten for the light, to cause the lamp to burn always.

Israhel ut adferant tibi oleum de arboribus oliuarum purissi-

mum piloque contusum ut ardeat lucerna semper 27:21 in 27:21 In the tabernacle of the congregation without the vail, which is before the testimony, Aaron and his sons shall order it from evening to morning before the LORD: it shall be a statute for ever unto their generations on the behalf of the children of Israel.

tabernaculo testimonii extra uelum quod oppansum est testimonio et conlocabunt eam Aaron et filii eius ut usque mane luceat coram Domino perpetuus erit cultus per successiones eorum a filiis Israhel

28:1 adplica quoque ad te Aaron fratrem tuum cum filiis 28:1 And take thou unto thee Aaron thy brother, and his

sons with him, from among the children of Israel, that he may minister unto me in the priest’s office, even Aaron,

suis de medio filiorum Israhel ut sacerdotio fungantur mihi Nadab and Abihu, Eleazar and Ithamar, Aaron’s sons. Aaron Nadab et Abiu Eleazar et Ithamar 28:2 faciesque ue- 28:2 And thou shalt make holy garments for Aaron thy

brother for glory and for beauty.

Biblia Sacra Vulgata

28:3 And thou shalt speak unto all that are wise hearted, stem sanctam fratri tuo in gloriam et decorem 28:3 et loque-

whom I have filled with the spirit of wisdom, that they may make Aaron’s garments to consecrate him, that he

may minister unto me in the priest’s office. ris cunctis sapientibus corde quos repleui spiritu prudentiae

ut faciant uestes Aaron in quibus sanctificatus ministret mihi

28:4 And these are the garments which they shall make; 28:4 haec autem erunt uestimenta quae facient rationale et

a breastplate, and an ephod, and a robe, and a broidered coat, a mitre, and a girdle: and they shall make holy garments for Aaron thy brother, and his sons, that he may minister unto me in the priest’s office.

superumerale tunicam et lineam strictam cidarim et balteum facient uestimenta sancta Aaron fratri tuo et filiis eius ut sa-

28:5 And they shall take gold, and blue, and purple, and cerdotio fungantur mihi 28:5 accipientque aurum et hyacin-

scarlet, and fine linen. 28:6 And they shall make the ephod of gold, of blue, and thum et purpuram coccumque bis tinctum et byssum 28:6 of purple, of scarlet, and fine twined linen, with cunning work.

facient autem superumerale de auro et hyacintho ac purpura

28:7 It shall have the two shoulderpieces thereof joined at coccoque bis tincto et bysso retorta opere polymito 28:7 the two edges thereof; and so it shall be joined together.

duas oras iunctas habebit in utroque latere summitatum ut in

28:8 And the curious girdle of the ephod, which is upon it, unum redeent 28:8 ipsaque textura et cuncta operis uarie-

shall be of the same, according to the work thereof; even of gold, of blue, and purple, and scarlet, and fine twined

linen. tas erit ex auro et hyacintho et purpura coccoque bis tincto 28:9 And thou shalt take two onyx stones, and grave on et bysso retorta 28:9 sumesque duos lapides onychinos et

them the names of the children of Israel:

28:10 Six of their names on one stone, and the other six sculpes in eis nomina filiorum Israhel 28:10 sex nomina in

names of the rest on the other stone, according to their birth.

lapide uno et sex reliqua in altero iuxta ordinem natiuitatis

28:11 With the work of an engraver in stone, like the en- eorum 28:11 opere sculptoris et celatura gemmarii sculpes

gravings of a signet, shalt thou engrave the two stones with the names of the children of Israel: thou shalt make them

to be set in ouches of gold. eos nominibus filiorum Israhel inclusos auro atque circum-28:12 And thou shalt put the two stones upon the shoulders datos 28:12 et pones in utroque latere superumeralis memo-

of the ephod for stones of memorial unto the children of Israel: and Aaron shall bear their names before the LORD

upon his two shoulders for a memorial. riale filiis Israhel portabitque Aaron nomina eorum coram 28:13 And thou shalt make ouches of gold; Domino super utrumque umerum ob recordationem 28:13 28:14 And two chains of pure gold at the ends; of wreathen facies et uncinos ex auro 28:14 et duas catenulas auri pu-

work shalt thou make them, and fasten the wreathen chains to the ouches.

rissimi sibi inuicem coherentes quas inseres uncinis 28:15

28:15 And thou shalt make the breastplate of judgment with cunning work; after the work of the ephod thou shalt make it; of gold, of blue, and of purple, and of scarlet, and of fine twined linen, shalt thou make it.

rationale quoque iudicii facies opere polymito iuxta texturam superumeralis ex auro hyacintho et purpura coccoque

28:16 Foursquare it shall be being doubled; a span shall be bis tincto et bysso retorta 28:16 quadrangulum erit et duthe length thereof, and a span shall be the breadth thereof.

plex mensuram palmi habebit tam in longitudine quam in

28:17 And thou shalt set in it settings of stones, even four latitudine 28:17 ponesque in eo quattuor ordines lapidum rows of stones: the first row shall be a sardius, a topaz, and a carbuncle: this shall be the first row.

in primo uersu erit lapis sardius et topazius et zmaragdus

28:18 And the second row shall be an emerald, a sapphire, 28:18 in secundo carbunculus sapphyrus et iaspis 28:19 in

and a diamond. 28:19 And the third row a ligure, an agate, and an amethyst.

tertio ligyrius achates et amethistus 28:20 in quarto chry-

28:20 And the fourth row a beryl, and an onyx, and a

jasper: they shall be set in gold in their inclosings. solitus onychinus et berillus inclusi auro erunt per ordines 28:21 And the stones shall be with the names of the chil- suos 28:21 habebuntque nomina filiorum Israhel duodecim

dren of Israel, twelve, according to their names, like the engravings of a signet; every one with his name shall they be according to the twelve tribes.

nominibus celabuntur singuli lapides nominibus singulorum per duodecim tribus 28:22 facies in rationali catenas sibi 28:22 And thou shalt make upon the breastplate chains at

the ends of wreathen work of pure gold.

inuicem coherentes ex auro purissimo 28:23 et duos anu- 28:23 And thou shalt make upon the breastplate two rings

of gold, and shalt put the two rings on the two ends of the breastplate.

los aureos quos pones in utraque rationalis summitate 28:24

28:24 And thou shalt put the two wreathen chains of gold in the two rings which are on the ends of the breastplate.

catenasque aureas iunges anulis qui sunt in marginibus eius 28:25 et ipsarum catenarum extrema duobus copulabis unci- 28:25 And the other two ends of the two wreathen chains

thou shalt fasten in the two ouches, and put them on the shoulderpieces of the ephod before it.

nis in utroque latere superumeralis quod rationale respicit 28:26 facies et duos anulos aureos quos pones in summitati- 28:26 And thou shalt make two rings of gold, and thou

shalt put them upon the two ends of the breastplate in the border thereof, which is in the side of the ephod inward.

bus rationalis et in oris quae e regione sunt superumeralis et posteriora eius aspiciunt 28:27 nec non et alios duos anu- 28:27 And two other rings of gold thou shalt make, and

shalt put them on the two sides of the ephod underneath, toward the forepart thereof, over against the other coupling

los aureos qui ponendi sunt in utroque latere superumera- thereof, above the curious girdle of the ephod. lis deorsum quod respicit contra faciem iuncturae inferioris ut aptari possit cum superumerali 28:28 et stringatur ratio- 28:28 And they shall bind the breastplate by the rings

thereof unto the rings of the ephod with a lace of blue, that it may be above the curious girdle of the ephod, and

nale anulis suis cum anulis superumeralis uitta hyacinthina that the breastplate be not loosed from the ephod. ut maneat iunctura fabrefacta et a se inuicem rationale et superumerale nequeant separari 28:29 portabitque Aaron no- 28:29 And Aaron shall bear the names of the children of

Israel in the breastplate of judgment upon his heart, when he goeth in unto the holy place, for a memorial before the

mina filiorum Israhel in rationali iudicii super pectus suum LORD continually. quando ingreditur sanctuarium memoriale coram Domino in aeternum 28:30 pones autem in rationali iudicii doctrinam 28:30 And thou shalt put in the breastplate of judgment the

Urim and the Thummim; and they shall be upon Aaron’s heart, when he goeth in before the LORD: and Aaron shall bear the judgment of the children of Israel upon his heart before the LORD continually.

et ueritatem quae erunt in pectore Aaron quando ingreditur coram Domino et gestabit iudicium filiorum Israhel in pectore suo in conspectu Domini semper 28:31 facies et tuni- 28:31 And thou shalt make the robe of the ephod all of

blue.

cam superumeralis totam hyacinthinam 28:32 in cuius me- 28:32 And there shall be an hole in the top of it, in the

midst thereof: it shall have a binding of woven work round about the hole of it, as it were the hole of an habergeon,

dio supra erit capitium et ora per gyrum eius textilis sicut that it be not rent. fieri solet in extremis uestium partibus ne facile rumpatur 28:33 deorsum uero ad pedes eiusdem tunicae per circuitum 28:33 And beneath upon the hem of it thou shalt make

pomegranates of blue, and of purple, and of scarlet, round about the hem thereof; and bells of gold between them

quasi mala punica facies ex hyacintho et purpura et cocco bis round about: tincto mixtis in medio tintinabulis 28:34 ita ut tintinabulum 28:34 A golden bell and a pomegranate, a golden bell and

a pomegranate, upon the hem of the robe round about.

sit aureum et malum rursumque tintinabulum aliud aureum et malum punicum 28:35 et uestietur ea Aaron in officio 28:35 And it shall be upon Aaron to minister: and his

sound shall be heard when he goeth in unto the holy place before the LORD, and when he cometh out, that he die

ministerii ut audiatur sonitus quando ingreditur et egreditur not. sanctuarium in conspectu Domini et non moriatur 28:36 fa- 28:36 And thou shalt make a plate of pure gold, and grave

upon it, like the engravings of a signet, HOLINESS TO THE LORD.

cies et lamminam de auro purissimo in qua sculpes opere

Biblia Sacra Vulgata

28:37 And thou shalt put it on a blue lace, that it may be celatoris Sanctum Domino 28:37 ligabisque eam uitta hya-

upon the mitre; upon the forefront of the mitre it shall be.

28:38 And it shall be upon Aaron’s forehead, that Aaron cinthina et erit super tiaram 28:38 inminens fronti pontificis

may bear the iniquity of the holy things, which the children of Israel shall hallow in all their holy gifts; and it shall be always upon his forehead, that they may be accepted before the LORD.

portabitque Aaron iniquitates eorum quae obtulerint et sanctificauerint filii Israhel in cunctis muneribus et donariis suis erit autem lammina semper in fronte eius ut placatus eis sit

28:39 And thou shalt embroider the coat of fine linen, and Dominus 28:39 stringesque tunicam bysso et tiaram byssi-

thou shalt make the mitre of fine linen, and thou shalt make the girdle of needlework.

nam facies et balteum opere plumarii 28:40 porro filiis Aa-

28:40 And for Aaron’s sons thou shalt make coats, and thou shalt make for them girdles, and bonnets shalt thou make for them, for glory and for beauty.

ron tunicas lineas parabis et balteos ac tiaras in gloriam et

28:41 And thou shalt put them upon Aaron thy brother, decorem 28:41 uestiesque his omnibus Aaron fratrem tuum

and his sons with him; and shalt anoint them, and consecrate them, and sanctify them, that they may minister unto

me in the priest’s office. et filios eius cum eo et cunctorum consecrabis manus sancti-28:42 And thou shalt make them linen breeches to cover ficabisque illos ut sacerdotio fungantur mihi 28:42 facies et

their nakedness; from the loins even unto the thighs they shall reach:

feminalia linea ut operiant carnem turpitudinis suae a reni-

28:43 And they shall be upon Aaron, and upon his sons, bus usque ad femina 28:43 et utentur eis Aaron et filii eius

when they come in unto the tabernacle of the congregation, or when they come near unto the altar to minister in the holy place; that they bear not iniquity, and die: it shall be a statute for ever unto him and his seed after him.

quando ingredientur tabernaculum testimonii uel quando adpropinquant ad altare ut ministrent in sanctuario ne iniquitatis rei moriantur legitimum sempiternum erit Aaron et semini eius post eum

29:1 And this is the thing that thou shalt do unto them 29:1 sed et hoc facies ut mihi in sacerdotio consecrentur

to hallow them, to minister unto me in the priest’s office: Take one young bullock, and two rams without blemish,

tolle uitulum de armento et arietes duos inmaculatos 29:2

29:2 And unleavened bread, and cakes unleavened tempered with oil, and wafers unleavened anointed with oil: of wheaten flour shalt thou make them.

panesque azymos et crustula absque fermento quae conspersa sint oleo lagana quoque azyma oleo lita de simila tri-

29:3 And thou shalt put them into one basket, and bring ticea cuncta facies 29:3 et posita in canistro offeres uitulum

them in the basket, with the bullock and the two rams.

29:4 And Aaron and his sons thou shalt bring unto the door autem et duos arietes 29:4 et Aaron ac filios eius adplicabis

of the tabernacle of the congregation, and shalt wash them with water.

ad ostium tabernaculi testimonii cumque laueris patrem cum

29:5 And thou shalt take the garments, and put upon Aaron filiis aqua 29:5 indues Aaron uestimentis suis id est linea et

the coat, and the robe of the ephod, and the ephod, and the breastplate, and gird him with the curious girdle of the

ephod: tunica et superumerali et rationali quod constringes balteo 29:6 And thou shalt put the mitre upon his head, and put 29:6 et pones tiaram in capite eius et lamminam sanctam su-

the holy crown upon the mitre.

29:7 Then shalt thou take the anointing oil, and pour it per tiaram 29:7 et oleum unctionis fundes super caput eius

upon his head, and anoint him.

29:8 And thou shalt bring his sons, and put coats upon atque hoc ritu consecrabitur 29:8 filios quoque illius adpli-

them.

29:9 And thou shalt gird them with girdles, Aaron and his cabis et indues tunicis lineis cingesque balteo 29:9 Aaron

sons, and put the bonnets on them: and the priest’s office shall be theirs for a perpetual statute: and thou shalt con-

secrate Aaron and his sons. scilicet et liberos eius et inpones eis mitras eruntque sacer-

dotes mei in religione perpetua postquam initiaueris manus

29:10 And thou shalt cause a bullock to be brought before eorum 29:10 adplicabis et uitulum coram tabernaculo tes-

the tabernacle of the congregation: and Aaron and his sons shall put their hands upon the head of the bullock.

timonii inponentque Aaron et filii eius manus super caput illius 29:11 et mactabis eum in conspectu Domini iuxta os- 29:11 And thou shalt kill the bullock before the LORD, by

the door of the tabernacle of the congregation.

tium tabernaculi testimonii 29:12 sumptumque de sanguine 29:12 And thou shalt take of the blood of the bullock, and

put it upon the horns of the altar with thy finger, and pour all the blood beside the bottom of the altar.

uituli pones super cornua altaris digito tuo reliquum autem

sanguinem fundes iuxta basim eius 29:13 sumes et adipem 29:13 And thou shalt take all the fat that covereth the inwards, and the caul that is above the liver, and the two kidneys, and the fat that is upon them, and burn them upon the

totum qui operit intestina et reticulum iecoris ac duos re- altar. nes et adipem qui super eos est et offeres incensum super altare 29:14 carnes uero uituli et corium et fimum conbu- 29:14 But the flesh of the bullock, and his skin, and his

dung, shalt thou burn with fire without the camp: it is a sin offering.

res foris extra castra eo quod pro peccato sit 29:15 unum

29:15 Thou shalt also take one ram; and Aaron and his sons shall put their hands upon the head of the ram.

quoque arietum sumes super cuius caput ponent Aaron et filii eius manus 29:16 quem cum mactaueris tolles de san- 29:16 And thou shalt slay the ram, and thou shalt take his

blood, and sprinkle it round about upon the altar.

guine eius et fundes circa altare 29:17 ipsum autem arie- 29:17 And thou shalt cut the ram in pieces, and wash the

inwards of him, and his legs, and put them unto his pieces, and unto his head.

tem secabis in frusta lotaque intestina eius ac pedes pones super concisas carnes et super caput illius 29:18 et offeres 29:18 And thou shalt burn the whole ram upon the altar: it

is a burnt offering unto the LORD: it is a sweet savour, an offering made by fire unto the LORD.

totum arietem in incensum super altare oblatio est Domini odor suauissimus uictimae Dei 29:19 tolles quoque arietem 29:19 And thou shalt take the other ram; and Aaron and

his sons shall put their hands upon the head of the ram.

alterum super cuius caput Aaron et filii eius ponent manus

29:20 quem cum immolaueris sumes de sanguine ipsius et 29:20 Then shalt thou kill the ram, and take of his blood, and put it upon the tip of the right ear of Aaron, and upon the tip of the right ear of his sons, and upon the thumb of their right hand, and upon the great toe of their right foot, and sprinkle the blood upon the altar round about.

pones super extremum dextrae auriculae Aaron et filiorum eius et super pollices manus eorum et pedis dextri fundesque sanguinem super altare per circuitum 29:21 cumque tuleris 29:21 And thou shalt take of the blood that is upon the

altar, and of the anointing oil, and sprinkle it upon Aaron, and upon his garments, and upon his sons, and upon the garments of his sons with him: and he shall be hallowed, and his garments, and his sons, and his sons’ garments with him.

de sanguine qui est super altare et de oleo unctionis asperges Aaron et uestes eius filios et uestimenta eorum consecratisque et ipsis et uestibus 29:22 tolles adipem de ariete et 29:22 Also thou shalt take of the ram the fat and the rump,

and the fat that covereth the inwards, and the caul above the liver, and the two kidneys, and the fat that is upon them, and the right shoulder; for it is a ram of consecration:

caudam et aruinam quae operit uitalia ac reticulum iecoris et duos renes atque adipem qui super eos est armumque dextrum eo quod sit aries consecrationum 29:23 tortam panis 29:23 And one loaf of bread, and one cake of oiled bread,

and one wafer out of the basket of the unleavened bread that is before the LORD:

unius crustulum conspersum oleo laganum de canistro azymorum quod positum est in conspectu Domini 29:24 pones- 29:24 And thou shalt put all in the hands of Aaron, and

in the hands of his sons; and shalt wave them for a wave offering before the LORD.

que omnia super manus Aaron et filiorum eius et sanctificabis eos eleuans coram Domino 29:25 suscipiesque uniuersa 29:25 And thou shalt receive them of their hands, and burn

them upon the altar for a burnt offering, for a sweet savour before the LORD: it is an offering made by fire unto the

de manibus eorum et incendes super altare in holocaustum LORD. odorem suauissimum in conspectu Domini quia oblatio eius

Biblia Sacra Vulgata

29:26 And thou shalt take the breast of the ram of Aaron’s est 29:26 sumes quoque pectusculum de ariete quo initia-

consecration, and wave it for a wave offering before the LORD: and it shall be thy part.

tus est Aaron sanctificabisque illud elatum coram Domino

29:27 And thou shalt sanctify the breast of the wave of- et cedet in partem tuam 29:27 sanctificabis et pectusculum

fering, and the shoulder of the heave offering, which is waved, and which is heaved up, of the ram of the consecration, even of that which is for Aaron, and of that which is for his sons:

consecratum et armum quem de ariete separasti 29:28 quo

29:28 And it shall be Aaron’s and his sons’ by a statute for ever from the children of Israel: for it is an heave offering: and it shall be an heave offering from the children of Israel of the sacrifice of their peace offerings, even their heave offering unto the LORD.

initiatus est Aaron et filii eius cedentque in partem Aaron et filiorum eius iure perpetuo a filiis Israhel quia primitiua sunt et initia de uictimis eorum pacificis quae offerunt Do-

29:29 And the holy garments of Aaron shall be his sons’ mino 29:29 uestem autem sanctam qua utitur Aaron habe-

after him, to be anointed therein, and to be consecrated in them.

bunt filii eius post eum ut unguantur in ea et consecrentur

29:30 And that son that is priest in his stead shall put them manus eorum 29:30 septem diebus utetur illa qui pontifex on seven days, when he cometh into the tabernacle of the congregation to minister in the holy place.

pro eo fuerit constitutus de filiis eius et qui ingredietur taber-

29:31 And thou shalt take the ram of the consecration, and naculum testimonii ut ministret in sanctuario 29:31 arietem

seethe his flesh in the holy place.

autem consecrationum tolles et coques carnes eius in loco

29:32 And Aaron and his sons shall eat the flesh of the sancto 29:32 quibus uescetur Aaron et filii eius panes quo-

ram, and the bread that is in the basket by the door of the tabernacle of the congregation.

que qui sunt in canistro in uestibulo tabernaculi testimonii

29:33 And they shall eat those things wherewith the atone- comedent 29:33 ut sit placabile sacrificium et sanctificen-

ment was made, to consecrate and to sanctify them: but a stranger shall not eat thereof, because they are holy.

tur offerentium manus alienigena non uescetur ex eis quia

29:34 And if ought of the flesh of the consecrations, or of sancti sunt 29:34 quod si remanserit de carnibus consecra-

the bread, remain unto the morning, then thou shalt burn the remainder with fire: it shall not be eaten, because it is

holy. tis siue de panibus usque mane conbures reliquias igni non 29:35 And thus shalt thou do unto Aaron, and to his sons, comedentur quia sanctificata sunt 29:35 omnia quae prae-

according to all things which I have commanded thee: seven days shalt thou consecrate them.

cepi tibi facies super Aaron et filiis eius septem diebus con-

29:36 And thou shalt offer every day a bullock for a sin secrabis manus eorum 29:36 et uitulum pro peccato offe-

offering for atonement: and thou shalt cleanse the altar, when thou hast made an atonement for it, and thou shalt

anoint it, to sanctify it. res per singulos dies ad expiandum mundabisque altare cum

immolaris expiationis hostiam et ungues illud in sanctifica-

29:37 Seven days thou shalt make an atonement for the tionem 29:37 septem diebus expiabis altare et sanctificabis

altar, and sanctify it; and it shall be an altar most holy: whatsoever toucheth the altar shall be holy.

et erit sanctum sanctorum omnis qui tetigerit illud sanctifi-

29:38 Now this is that which thou shalt offer upon the al- cabitur 29:38 hoc est quod facies in altari agnos anniculos

tar; two lambs of the first year day by day continually.

29:39 The one lamb thou shalt offer in the morning; and duos per singulos dies iugiter 29:39 unum agnum mane et

the other lamb thou shalt offer at even:

29:40 And with the one lamb a tenth deal of flour mingled alterum uespere 29:40 decimam partem similae conspersae

with the fourth part of an hin of beaten oil; and the fourth part of an hin of wine for a drink offering.

oleo tunso quod habeat mensuram quartam partem hin et ui-

29:41 And the other lamb thou shalt offer at even, and shalt num ad libandum eiusdem mensurae in agno uno 29:41 aldo thereto according to the meat offering of the morning, and according to the drink offering thereof, for a sweet

savour, an offering made by fire unto the LORD. terum uero agnum offeres ad uesperam iuxta ritum matutinae oblationis et iuxta ea quae diximus in odorem suauitatis 29:42 sacrificium Domino oblatione perpetua in generatio- 29:42 This shall be a continual burnt offering through-

out your generations at the door of the tabernacle of the congregation before the LORD: where I will meet you, to

nes uestras ad ostium tabernaculi testimonii coram Domino speak there unto thee. ubi constituam ut loquar ad te 29:43 ibique praecipiam filiis 29:43 And there I will meet with the children of Israel, and

the tabernacle shall be sanctified by my glory.

Israhel et sanctificabitur altare in gloria mea 29:44 sancti- 29:44 And I will sanctify the tabernacle of the congrega-

tion, and the altar: I will sanctify also both Aaron and his sons, to minister to me in the priest’s office.

ficabo et tabernaculum testimonii cum altari et Aaron cum filiis eius ut sacerdotio fungantur mihi 29:45 et habitabo in 29:45 And I will dwell among the children of Israel, and

will be their God.

medio filiorum Israhel eroque eis Deus 29:46 et scient quia 29:46 And they shall know that I am the LORD their God,

that brought them forth out of the land of Egypt, that I may dwell among them: I am the LORD their God.

ego Dominus Deus eorum qui eduxi eos de terra Aegypti ut manerem inter illos ego Dominus Deus ipsorum

30:1 facies quoque altare in adolendum thymiama de li- 30:1 And thou shalt make an altar to burn incense upon:

of shittim wood shalt thou make it.

gnis setthim 30:2 habens cubitum longitudinis et alterum 30:2 A cubit shall be the length thereof, and a cubit the

breadth thereof; foursquare shall it be: and two cubits shall be the height thereof: the horns thereof shall be of the

latitudinis id est quadrangulum et duos cubitos in altitudine same. cornua ex ipso procedent 30:3 uestiesque illud auro puris- 30:3 And thou shalt overlay it with pure gold, the top

thereof, and the sides thereof round about, and the horns thereof; and thou shalt make unto it a crown of gold round

simo tam craticulam eius quam parietes per circuitum et cor- about. nua faciesque ei coronam aureolam per gyrum 30:4 et duos 30:4 And two golden rings shalt thou make to it under the

crown of it, by the two corners thereof, upon the two sides of it shalt thou make it; and they shall be for places for the

anulos aureos sub corona per singula latera ut mittantur in staves to bear it withal. eos uectes et altare portetur 30:5 ipsos quoque uectes facies 30:5 And thou shalt make the staves of shittim wood, and

overlay them with gold.

de lignis setthim et inaurabis 30:6 ponesque altare contra 30:6 And thou shalt put it before the vail that is by the

ark of the testimony, before the mercy seat that is over the testimony, where I will meet with thee.

uelum quod ante arcam pendet testimonii coram propitiatorio quo tegitur testimonium ubi loquar tibi 30:7 et adolebit 30:7 And Aaron shall burn thereon sweet incense every

morning: when he dresseth the lamps, he shall burn incense upon it.

incensum super eo Aaron suaue fraglans mane quando conponet lucernas incendet illud 30:8 et quando conlocat eas 30:8 And when Aaron lighteth the lamps at even, he shall

burn incense upon it, a perpetual incense before the LORD throughout your generations.

ad uesperum uret thymiama sempiternum coram Domino in generationes uestras 30:9 non offeretis super eo thymiama 30:9 Ye shall offer no strange incense thereon, nor burnt

sacrifice, nor meat offering; neither shall ye pour drink offering thereon.

conpositionis alterius nec oblationem et uictimam nec liba libabitis 30:10 et deprecabitur Aaron super cornua eius se- 30:10 And Aaron shall make an atonement upon the horns

of it once in a year with the blood of the sin offering of atonements: once in the year shall he make atonement upon it throughout your generations: it is most holy unto the LORD.

mel per annum in sanguine quod oblatum est pro peccato et placabit super eo in generationibus uestris sanctum sanctorum erit Domino 30:11 locutusque est Dominus ad Mosen 30:11 And the LORD spake unto Moses, saying, dicens 30:12 quando tuleris summam filiorum Israhel iuxta 30:12 When thou takest the sum of the children of Israel

after their number, then shall they give every man a ransom for his soul unto the LORD, when thou numberest them; that there be no plague among them, when thou numberest them.

numerum dabunt singuli pretium pro animabus suis Domino et non erit plaga in eis cum fuerint recensiti 30:13 hoc au- 30:13 This they shall give, every one that passeth among

them that are numbered, half a shekel after the shekel of the sanctuary: (a shekel is twenty gerahs:) an half shekel

tem dabit omnis qui transit ad nomen dimidium sicli iuxta shall be the offering of the LORD.

mensuram templi siclus uiginti obolos habet media pars sicli

30:14 Every one that passeth among them that are num- offeretur Domino 30:14 qui habetur in numero a uiginti an-

bered, from twenty years old and above, shall give an offering unto the LORD.

nis et supra dabit pretium 30:15 diues non addet ad medium

30:15 The rich shall not give more, and the poor shall not give less than half a shekel, when they give an offering unto the LORD, to make an atonement for your souls.

sicli et pauper nihil minuet 30:16 susceptamque pecuniam

30:16 And thou shalt take the atonement money of the children of Israel, and shalt appoint it for the service of the tabernacle of the congregation; that it may be a memorial unto the children of Israel before the LORD, to make an atonement for your souls.

quae conlata est a filiis Israhel trades in usus tabernaculi testimonii ut sit monumentum eorum coram Domino et pro-

30:17 And the LORD spake unto Moses, saying, pitietur animabus illorum 30:17 locutusque est Dominus ad

30:18 Thou shalt also make a laver of brass, and his foot Mosen dicens 30:18 facies et labium aeneum cum basi sua also of brass, to wash withal: and thou shalt put it between the tabernacle of the congregation and the altar, and thou

shalt put water therein. ad lauandum ponesque illud inter tabernaculum testimonii et 30:19 For Aaron and his sons shall wash their hands and altare et missa aqua 30:19 lauabunt in eo Aaron et filii eius

their feet thereat:

30:20 When they go into the tabernacle of the congrega- manus suas ac pedes 30:20 quando ingressuri sunt taberna-

tion, they shall wash with water, that they die not; or when they come near to the altar to minister, to burn offering

made by fire unto the LORD: culum testimonii et quando accessuri ad altare ut offerant in

30:21 So they shall wash their hands and their feet, that eo thymiama Domino 30:21 ne forte moriantur legitimum they die not: and it shall be a statute for ever to them, even to him and to his seed throughout their generations.

sempiternum erit ipsi et semini eius per successiones 30:22

30:22 Moreover the LORD spake unto Moses, saying,

30:23 Take thou also unto thee principal spices, of pure locutusque est Dominus ad Mosen 30:23 dicens sume tibi myrrh five hundred shekels, and of sweet cinnamon half so much, even two hundred and fifty shekels, and of sweet

calamus two hundred and fifty shekels, aromata prima et zmyrnae electae quingentos siclos et cinnamomi medium id est ducentos quinquaginta calami simili-

30:24 And of cassia five hundred shekels, after the shekel ter ducentos quinquaginta 30:24 cassiae autem quingentos

of the sanctuary, and of oil olive an hin:

siclos in pondere sanctuarii olei de oliuetis mensuram hin

30:25 And thou shalt make it an oil of holy ointment, an 30:25 faciesque unctionis oleum sanctum unguentum conpo-

ointment compound after the art of the apothecary: it shall be an holy anointing oil.

situm opere unguentarii 30:26 et ungues ex eo tabernacu-

30:26 And thou shalt anoint the tabernacle of the congregation therewith, and the ark of the testimony,

30:27 And the table and all his vessels, and the candlestick lum testimonii et arcam testamenti 30:27 mensamque cum and his vessels, and the altar of incense,

uasis suis candelabrum et utensilia eius altaria thymiamatis

30:28 And the altar of burnt offering with all his vessels, 30:28 et holocausti et uniuersam supellectilem quae ad cul-

and the laver and his foot.

30:29 And thou shalt sanctify them, that they may be most tum eorum pertinent 30:29 sanctificabisque omnia et erunt

holy: whatsoever toucheth them shall be holy.

30:30 And thou shalt anoint Aaron and his sons, and con- sancta sanctorum qui tetigerit ea sanctificabitur 30:30 Aa-

secrate them, that they may minister unto me in the priest’s office.

ron et filios eius ungues sanctificabisque eos ut sacerdotio

30:31 And thou shalt speak unto the children of Israel, say- fungantur mihi 30:31 filiis quoque Israhel dices hoc oleum ing, This shall be an holy anointing oil unto me throughout your generations.

unctionis sanctum erit mihi in generationes uestras 30:32

30:32 Upon man’s flesh shall it not be poured, neither shall ye make any other like it, after the composition of it: it is holy, and it shall be holy unto you.

caro hominis non unguetur ex eo et iuxta conpositionem eius non facietis aliud quia sanctificatum est et sanctum erit uo-

30:33 Whosoever compoundeth any like it, or whosoever bis 30:33 homo quicumque tale conposuerit et dederit ex eo

putteth any of it upon a stranger, shall even be cut off from his people.

alieno exterminabitur de populo suo 30:34 dixitque Domi-

30:34 And the LORD said unto Moses, Take unto thee sweet spices, stacte, and onycha, and galbanum; these sweet spices with pure frankincense: of each shall there be a like weight:

nus ad Mosen sume tibi aromata stacten et onycha galbanen boni odoris et tus lucidissimum aequalis ponderis erunt om-

nia 30:35 faciesque thymiama conpositum opere unguen- 30:35 And thou shalt make it a perfume, a confection after

the art of the apothecary, tempered together, pure and holy:

tarii mixtum diligenter et purum et sanctificatione dignissimum 30:36 cumque in tenuissimum puluerem uniuersa 30:36 And thou shalt beat some of it very small, and put

of it before the testimony in the tabernacle of the congregation, where I will meet with thee: it shall be unto you

contuderis pones ex eo coram testimonio tabernaculi in quo most holy. loco apparebo tibi sanctum sanctorum erit uobis thymiama 30:37 talem conpositionem non facietis in usus uestros quia 30:37 And as for the perfume which thou shalt make, ye

shall not make to yourselves according to the composition thereof: it shall be unto thee holy for the LORD.

sanctum est Domino 30:38 homo quicumque fecerit simile

30:38 Whosoever shall make like unto that, to smell thereto, shall even be cut off from his people.

ut odore illius perfruatur peribit de populis suis

31:1 locutusque est Dominus ad Mosen dicens 31:2 ecce 31:1 And the LORD spake unto Moses, saying,

31:2 See, I have called by name Bezaleel the son of Uri, the son of Hur, of the tribe of Judah:

uocaui ex nomine Beselehel filium Vri filii Hur de tribu Iuda 31:3 et impleui eum spiritu Dei sapientia intellegentia 31:3 And I have filled him with the spirit of God, in wis-

dom, and in understanding, and in knowledge, and in all manner of workmanship,

et scientia in omni opere 31:4 ad excogitandum fabre quic-

31:4 To devise cunning works, to work in gold, and in silver, and in brass,

quid fieri potest ex auro et argento et aere 31:5 marmore 31:5 And in cutting of stones, to set them, and in carving

of timber, to work in all manner of workmanship.

et gemmis et diuersitate lignorum 31:6 dedique ei socium 31:6 And I, behold, I have given with him Aholiab, the

son of Ahisamach, of the tribe of Dan: and in the hearts of all that are wise hearted I have put wisdom, that they may

Hooliab filium Achisamech de tribu Dan et in corde omnis make all that I have commanded thee; eruditi posui sapientiam ut faciant cuncta quae praecepi tibi 31:7 tabernaculum foederis et arcam testimonii et propitia- 31:7 The tabernacle of the congregation, and the ark of the

testimony, and the mercy seat that is thereupon, and all the furniture of the tabernacle,

torium quod super eam est et cuncta uasa tabernaculi 31:8

31:8 And the table and his furniture, and the pure candlestick with all his furniture, and the altar of incense,

mensamque et uasa eius candelabrum purissimum cum uasis suis et altaria thymiamatis 31:9 et holocausti et omnia uasa 31:9 And the altar of burnt offering with all his furniture,

and the laver and his foot,

eorum labium cum basi sua 31:10 uestes sanctas in mini- 31:10 And the cloths of service, and the holy garments for Aaron the priest, and the garments of his sons, to minister in the priest’s office,

sterio Aaron sacerdoti et filiis eius ut fungantur officio suo

in sacris 31:11 oleum unctionis et thymiama aromatum in 31:11 And the anointing oil, and sweet incense for the holy place: according to all that I have commanded thee shall they do.

sanctuario omnia quae praecepi tibi facient 31:12 et locu-

31:12 And the LORD spake unto Moses, saying,

tus est Dominus ad Mosen dicens 31:13 loquere filiis Is- 31:13 Speak thou also unto the children of Israel, say-

ing, Verily my sabbaths ye shall keep: for it is a sign between me and you throughout your generations; that ye

rahel et dices ad eos uidete ut sabbatum meum custodiatis may know that I am the LORD that doth sanctify you. quia signum est inter me et uos in generationibus uestris ut sciatis quia ego Dominus qui sanctifico uos 31:14 custodite 31:14 Ye shall keep the sabbath therefore; for it is holy

unto you: every one that defileth it shall surely be put to death: for whosoever doeth any work therein, that soul

sabbatum sanctum est enim uobis qui polluerit illud morte shall be cut off from among his people. morietur qui fecerit in eo opus peribit anima illius de medio populi sui 31:15 sex diebus facietis opus in die septimo 31:15 Six days may work be done; but in the seventh is the

sabbath of rest, holy to the LORD: whosoever doeth any work in the sabbath day, he shall surely be put to death.

Biblia Sacra Vulgata

sabbatum est requies sancta Domino omnis qui fecerit opus

31:16 Wherefore the children of Israel shall keep the sab- in hac die morietur 31:16 custodiant filii Israhel sabbatum

bath, to observe the sabbath throughout their generations, for a perpetual covenant.

et celebrent illud in generationibus suis pactum est sempiter-

31:17 It is a sign between me and the children of Israel for num 31:17 inter me et filios Israhel signumque perpetuum ever: for in six days the LORD made heaven and earth, and on the seventh day he rested, and was refreshed.

sex enim diebus fecit Dominus caelum et terram et in sep-

31:18 And he gave unto Moses, when he had made an end timo ab opere cessauit 31:18 dedit quoque Mosi conpletis

of communing with him upon mount Sinai, two tables of testimony, tables of stone, written with the finger of God.

huiuscemodi sermonibus in monte Sinai duas tabulas testimonii lapideas scriptas digito Dei

32:1 And when the people saw that Moses delayed to 32:1 uidens autem populus quod moram faceret descen-

come down out of the mount, the people gathered themselves together unto Aaron, and said unto him, Up, make us gods, which shall go before us; for as for this Moses, the man that brought us up out of the land of Egypt, we wot not what is become of him.

dendi de monte Moses congregatus aduersus Aaron ait surge fac nobis deos qui nos praecedant Mosi enim huic uiro qui

32:2 And Aaron said unto them, Break off the golden ear- nos eduxit de terra Aegypti ignoramus quid acciderit 32:2

rings, which are in the ears of your wives, of your sons, and of your daughters, and bring them unto me.

dixitque ad eos Aaron tollite inaures aureas de uxorum filio-

32:3 And all the people brake off the golden earrings rumque et filiarum uestrarum auribus et adferte ad me 32:3

which were in their ears, and brought them unto Aaron.

32:4 And he received them at their hand, and fashioned it fecit populus quae iusserat deferens inaures ad Aaron 32:4

with a graving tool, after he had made it a molten calf: and they said, These be thy gods, O Israel, which brought thee

up out of the land of Egypt. quas cum ille accepisset formauit opere fusorio et fecit ex

eis uitulum conflatilem dixeruntque hii sunt dii tui Israhel

32:5 And when Aaron saw it, he built an altar before it; qui te eduxerunt de terra Aegypti 32:5 quod cum uidisset

and Aaron made proclamation, and said, To morrow is a feast to the LORD.

Aaron aedificauit altare coram eo et praeconis uoce clama-

32:6 And they rose up early on the morrow, and offered uit dicens cras sollemnitas Domini est 32:6 surgentesque

burnt offerings, and brought peace offerings; and the people sat down to eat and to drink, and rose up to play.

mane obtulerunt holocausta et hostias pacificas et sedit po-

32:7 And the LORD said unto Moses, Go, get thee down; pulus comedere ac bibere et surrexerunt ludere 32:7 locu-

for thy people, which thou broughtest out of the land of Egypt, have corrupted themselves:

tus est autem Dominus ad Mosen uade descende peccauit

32:8 They have turned aside quickly out of the way which populus tuus quem eduxisti de terra Aegypti 32:8 recesse-

I commanded them: they have made them a molten calf, and have worshipped it, and have sacrificed thereunto, and said, These be thy gods, O Israel, which have brought thee up out of the land of Egypt.

runt cito de uia quam ostendisti eis feceruntque sibi uitulum conflatilem et adorauerunt atque immolantes ei hostias dixerunt isti sunt dii tui Israhel qui te eduxerunt de terra Ae-

32:9 And the LORD said unto Moses, I have seen this gypti 32:9 rursumque ait Dominus ad Mosen cerno quod

people, and, behold, it is a stiffnecked people:

32:10 Now therefore let me alone, that my wrath may wax populus iste durae ceruicis sit 32:10 dimitte me ut irascatur

hot against them, and that I may consume them: and I will make of thee a great nation.

furor meus contra eos et deleam eos faciamque te in gen-

32:11 And Moses besought the LORD his God, and said, tem magnam 32:11 Moses autem orabat Dominum Deum LORD, why doth thy wrath wax hot against thy people, which thou hast brought forth out of the land of Egypt

with great power, and with a mighty hand? suum dicens cur Domine irascitur furor tuus contra populum tuum quem eduxisti de terra Aegypti in fortitudine magna et in manu robusta 32:12 ne quaeso dicant Aegyptii callide 32:12 Wherefore should the Egyptians speak, and say, For

mischief did he bring them out, to slay them in the mountains, and to consume them from the face of the earth? Turn from thy fierce wrath, and repent of this evil against thy people.

eduxit eos ut interficeret in montibus et deleret e terra quiescat ira tua et esto placabilis super nequitia populi tui 32:13 32:13 Remember Abraham, Isaac, and Israel, thy servants,

to whom thou swarest by thine own self, and saidst unto them, I will multiply your seed as the stars of heaven, and all this land that I have spoken of will I give unto your seed, and they shall inherit it for ever.

recordare Abraham Isaac et Israhel seruorum tuorum quibus iurasti per temet ipsum dicens multiplicabo semen uestrum sicut stellas caeli et uniuersam terram hanc de qua locutus sum dabo semini uestro et possidebitis eam semper 32:14 32:14 And the LORD repented of the evil which he

thought to do unto his people.

placatusque est Dominus ne faceret malum quod locutus fuerat aduersus populum suum 32:15 et reuersus est Moses de 32:15 And Moses turned, and went down from the mount,

and the two tables of the testimony were in his hand: the tables were written on both their sides; on the one side and

monte portans duas tabulas testimonii manu scriptas ex ut- on the other were they written. raque parte 32:16 et factas opere Dei scriptura quoque Dei 32:16 And the tables were the work of God, and the writ-

ing was the writing of God, graven upon the tables.

erat sculpta in tabulis 32:17 audiens autem Iosue tumultum 32:17 And when Joshua heard the noise of the people as

they shouted, he said unto Moses, There is a noise of war in the camp.

populi uociferantis dixit ad Mosen ululatus pugnae auditur in castris 32:18 qui respondit non est clamor adhortantium 32:18 And he said, It is not the voice of them that shout for

mastery, neither is it the voice of them that cry for being overcome: but the noise of them that sing do I hear.

ad pugnam neque uociferatio conpellentium ad fugam sed uocem cantantium ego audio 32:19 cumque adpropinquas- 32:19 And it came to pass, as soon as he came nigh unto

the camp, that he saw the calf, and the dancing: and Moses’ anger waxed hot, and he cast the tables out of his

set ad castra uidit uitulum et choros iratusque ualde proiecit hands, and brake them beneath the mount. de manu tabulas et confregit eas ad radices montis 32:20 32:20 And he took the calf which they had made, and burnt

it in the fire, and ground it to powder, and strawed it upon the water, and made the children of Israel drink of it.

arripiensque uitulum quem fecerant conbusit et contriuit usque ad puluerem quem sparsit in aqua et dedit ex eo potum filiis Israhel 32:21 dixitque ad Aaron quid tibi fecit hic 32:21 And Moses said unto Aaron, What did this people

unto thee, that thou hast brought so great a sin upon them?

populus ut induceres super eum peccatum maximum 32:22 32:22 And Aaron said, Let not the anger of my lord wax

hot: thou knowest the people, that they are set on mischief.

cui ille respondit ne indignetur dominus meus tu enim nosti populum istum quod pronus sit ad malum 32:23 dixerunt 32:23 For they said unto me, Make us gods, which shall

go before us: for as for this Moses, the man that brought us up out of the land of Egypt, we wot not what is become

mihi fac nobis deos qui praecedant nos huic enim Mosi qui of him. nos eduxit de terra Aegypti nescimus quid acciderit 32:24 32:24 And I said unto them, Whosoever hath any gold, let

them break it off. So they gave it me: then I cast it into the fire, and there came out this calf.

quibus ego dixi quis uestrum habet aurum tulerunt et dederunt mihi et proieci illud in ignem egressusque est hic uitulus 32:25 uidens ergo Moses populum quod esset nudatus spolia- 32:25 And when Moses saw that the people were naked;

(for Aaron had made them naked unto their shame among their enemies:)

uerat enim eum Aaron propter ignominiam sordis et inter hostes nudum constituerat 32:26 et stans in porta castrorum ait 32:26 Then Moses stood in the gate of the camp, and said,

Who is on the LORD’s side? let him come unto me. And all the sons of Levi gathered themselves together unto him.

si quis est Domini iungatur mihi congregatique sunt ad eum omnes filii Leui 32:27 quibus ait haec dicit Dominus De- 32:27 And he said unto them, Thus saith the LORD God

of Israel, Put every man his sword by his side, and go in and out from gate to gate throughout the camp, and slay every man his brother, and every man his companion, and every man his neighbour.

Biblia Sacra Vulgata

us Israhel ponat uir gladium super femur suum ite et rediie de porta usque ad portam per medium castrorum et occidat

32:28 And the children of Levi did according to the word unusquisque fratrem et amicum et proximum suum 32:28 of Moses: and there fell of the people that day about three thousand men.

fecerunt filii Leui iuxta sermonem Mosi cecideruntque in die

32:29 For Moses had said, Consecrate yourselves today illo quasi tria milia hominum 32:29 et ait Moses consecras-

to the LORD, even every man upon his son, and upon his brother; that he may bestow upon you a blessing this day.

tis manus uestras hodie Domino unusquisque in filio et fratre

32:30 And it came to pass on the morrow, that Moses said suo ut detur uobis benedictio 32:30 facto autem die altero

unto the people, Ye have sinned a great sin: and now I will go up unto the LORD; peradventure I shall make an

atonement for your sin. locutus est Moses ad populum peccastis peccatum maximum

ascendam ad Dominum si quo modo eum quiuero deprecari

32:31 And Moses returned unto the LORD, and said, Oh, pro scelere uestro 32:31 reuersusque ad Dominum ait ob-

this people have sinned a great sin, and have made them gods of gold.

secro peccauit populus iste peccatum magnum feceruntque

32:32 Yet now, if thou wilt forgive their sin–; and if not, sibi deos aureos aut dimitte eis hanc noxam 32:32 aut si non

blot me, I pray thee, out of thy book which thou hast written.

facis dele me de libro tuo quem scripsisti 32:33 cui respon-

32:33 And the LORD said unto Moses, Whosoever hath sinned against me, him will I blot out of my book.

dit Dominus qui peccauerit mihi delebo eum de libro meo

32:34 Therefore now go, lead the people unto the place of 32:34 tu autem uade et duc populum istum quo locutus sum which I have spoken unto thee: behold, mine Angel shall go before thee: nevertheless in the day when I visit I will

visit their sin upon them. tibi angelus meus praecedet te ego autem in die ultionis uisi-32:35 And the LORD plagued the people, because they tabo et hoc peccatum eorum 32:35 percussit ergo Dominus

made the calf, which Aaron made.

populum pro reatu uituli quem fecit Aaron

33:1 And the LORD said unto Moses, Depart, and go up 33:1 locutusque est Dominus ad Mosen uade ascende de

hence, thou and the people which thou hast brought up out of the land of Egypt, unto the land which I sware unto Abraham, to Isaac, and to Jacob, saying, Unto thy seed will I give it:

loco isto tu et populus tuus quem eduxisti de terra Aegypti in terram quam iuraui Abraham Isaac et Iacob dicens se-

33:2 And I will send an angel before thee; and I will drive mini tuo dabo eam 33:2 et mittam praecursorem tui anout the Canaanite, the Amorite, and the Hittite, and the Perizzite, the Hivite, and the Jebusite:

gelum ut eiciam Chananeum et Amorreum et Hettheum et

33:3 Unto a land flowing with milk and honey: for I will Ferezeum et Eueum et Iebuseum 33:3 et intres in terram

not go up in the midst of thee; for thou art a stiffnecked people: lest I consume thee in the way.

fluentem lacte et melle non enim ascendam tecum quia po-

33:4 And when the people heard these evil tidings, they pulus durae ceruicis est ne forte disperdam te in uia 33:4

mourned: and no man did put on him his ornaments.

audiens populus sermonem hunc pessimum luxit et nullus

33:5 For the LORD had said unto Moses, Say unto the ex more indutus est cultu suo 33:5 dixitque Dominus ad

children of Israel, Ye are a stiffnecked people: I will come up into the midst of thee in a moment, and consume thee: therefore now put off thy ornaments from thee, that I may know what to do unto thee.

Mosen loquere filiis Israhel populus durae ceruicis es semel ascendam in medio tui et delebo te iam nunc depone

33:6 And the children of Israel stripped themselves of their ornatum tuum ut sciam quid faciam tibi 33:6 deposuerunt ornaments by the mount Horeb. 33:7 And Moses took the tabernacle, and pitched it with- ergo filii Israhel ornatum suum a monte Horeb 33:7 Moout the camp, afar off from the camp, and called it the Tabernacle of the congregation. And it came to pass, that every one which sought the LORD went out unto the tabernacle of the congregation, which was without the camp.

ses quoque tollens tabernaculum tetendit extra castra procul uocauitque nomen eius tabernaculum foederis et omnis populus qui habebat aliquam quaestionem egrediebatur ad tabernaculum foederis extra castra 33:8 cumque egredere- 33:8 And it came to pass, when Moses went out unto the

tabernacle, that all the people rose up, and stood every man at his tent door, and looked after Moses, until he was gone

tur Moses ad tabernaculum surgebat uniuersa plebs et sta- into the tabernacle. bat unusquisque in ostio papilionis sui aspiciebantque tergum Mosi donec ingrederetur tentorium 33:9 ingresso au- 33:9 And it came to pass, as Moses entered into the taber-

nacle, the cloudy pillar descended, and stood at the door of the tabernacle, and the Lord talked with Moses.

tem illo tabernaculum foederis descendebat columna nubis et stabat ad ostium loquebaturque cum Mosi 33:10 cernen- 33:10 And all the people saw the cloudy pillar stand at

the tabernacle door: and all the people rose up and worshipped, every man in his tent door.

tibus uniuersis quod columna nubis staret ad ostium tabernaculi stabantque ipsi et adorabant per fores tabernaculorum suorum 33:11 loquebatur autem Dominus ad Mosen facie 33:11 And the LORD spake unto Moses face to face, as

a man speaketh unto his friend. And he turned again into the camp: but his servant Joshua, the son of Nun, a young

ad faciem sicut loqui solet homo ad amicum suum cumque man, departed not out of the tabernacle. ille reuerteretur in castra minister eius Iosue filius Nun puer non recedebat de tabernaculo 33:12 dixit autem Moses ad 33:12 And Moses said unto the LORD, See, thou sayest

unto me, Bring up this people: and thou hast not let me know whom thou wilt send with me. Yet thou hast said, I know thee by name, and thou hast also found grace in my sight.

Dominum praecipis ut educam populum istum et non indicas mihi quem missurus es mecum praesertim cum dixeris noui

te ex nomine et inuenisti gratiam coram me 33:13 si ergo 33:13 Now therefore, I pray thee, if I have found grace in thy sight, shew me now thy way, that I may know thee, that I may find grace in thy sight: and consider that this

inueni gratiam in conspectu tuo ostende mihi uiam tuam ut nation is thy people. sciam te et inueniam gratiam ante oculos tuos respice populum tuum gentem hanc 33:14 dixitque Dominus facies mea 33:14 And he said, My presence shall go with thee, and I

will give thee rest.

praecedet te et requiem dabo tibi 33:15 et ait Moses si non 33:15 And he said unto him, If thy presence go not with

me, carry us not up hence.

tu ipse praecedes ne educas nos de loco isto 33:16 in quo 33:16 For wherein shall it be known here that I and thy

people have found grace in thy sight? is it not in that thou goest with us? so shall we be separated, I and thy people,

enim scire poterimus ego et populus tuus inuenisse nos gra- from all the people that are upon the face of the earth. tiam in conspectu tuo nisi ambulaueris nobiscum ut glorificemur ab omnibus populis qui habitant super terram 33:17 33:17 And the LORD said unto Moses, I will do this thing

also that thou hast spoken: for thou hast found grace in my sight, and I know thee by name.

dixit autem Dominus ad Mosen et uerbum istud quod locutus es faciam inuenisti enim gratiam coram me et te ipsum noui ex nomine 33:18 qui ait ostende mihi gloriam tuam 33:18 And he said, I beseech thee, shew me thy glory. 33:19 respondit ego ostendam omne bonum tibi et uocabo in 33:19 And he said, I will make all my goodness pass be-

fore thee, and I will proclaim the name of the LORD before thee; and will be gracious to whom I will be gracious,

nomine Domini coram te et miserebor cui uoluero et cle- and will shew mercy on whom I will shew mercy. mens ero in quem mihi placuerit 33:20 rursumque ait non 33:20 And he said, Thou canst not see my face: for there

shall no man see me, and live.

poteris uidere faciem meam non enim uidebit me homo et uiuet 33:21 et iterum ecce inquit est locus apud me stabis 33:21 And the LORD said, Behold, there is a place by me,

and thou shalt stand upon a rock:

Biblia Sacra Vulgata

33:22 And it shall come to pass, while my glory passeth super petram 33:22 cumque transibit gloria mea ponam te

by, that I will put thee in a clift of the rock, and will cover thee with my hand while I pass by:

in foramine petrae et protegam dextera mea donec transeam

33:23 And I will take away mine hand, and thou shalt see 33:23 tollamque manum meam et uidebis posteriora mea fa-

my back parts: but my face shall not be seen.

ciem autem meam uidere non poteris

34:1 And the LORD said unto Moses, Hew thee two tables 34:1 ac deinceps praecide ait tibi duas tabulas lapideas in-

of stone like unto the first: and I will write upon these tables the words that were in the first tables, which thou

brakest. star priorum et scribam super eas uerba quae habuerunt ta-34:2 And be ready in the morning, and come up in the bulae quas fregisti 34:2 esto paratus mane ut ascendas sta-

morning unto mount Sinai, and present thyself there to me in the top of the mount.

tim in montem Sinai stabisque mecum super uerticem mon-

34:3 And no man shall come up with thee, neither let tis 34:3 nullus ascendat tecum nec uideatur quispiam per

any man be seen throughout all the mount; neither let the flocks nor herds feed before that mount.

totum montem boues quoque et oues non pascantur e con-

34:4 And he hewed two tables of stone like unto the first; tra 34:4 excidit ergo duas tabulas lapideas quales ante fue-

and Moses rose up early in the morning, and went up unto mount Sinai, as the LORD had commanded him, and took

in his hand the two tables of stone. rant et de nocte consurgens ascendit in montem Sinai sicut 34:5 And the LORD descended in the cloud, and stood ei praeceperat Dominus portans secum tabulas 34:5 cum-

with him there, and proclaimed the name of the LORD.

que descendisset Dominus per nubem stetit Moses cum eo

34:6 And the LORD passed by before him, and pro- inuocans nomen Domini 34:6 quo transeunte coram eo ait

claimed, The LORD, The LORD God, merciful and gracious, longsuffering, and abundant in goodness and truth,

Dominator Domine Deus misericors et clemens patiens et

34:7 Keeping mercy for thousands, forgiving iniquity and multae miserationis ac uerus 34:7 qui custodis misericor-

transgression and sin, and that will by no means clear the guilty; visiting the iniquity of the fathers upon the children, and upon the children’s children, unto the third and to the fourth generation.

diam in milia qui aufers iniquitatem et scelera atque peccata nullusque apud te per se innocens est qui reddis iniquitatem patrum in filiis ac nepotibus in tertiam et quartam pro-

34:8 And Moses made haste, and bowed his head toward geniem 34:8 festinusque Moses curuatus est pronus in ter-

the earth, and worshipped.

34:9 And he said, If now I have found grace in thy sight, ram et adorans 34:9 ait si inueni gratiam in conspectu tuo

O LORD, let my LORD, I pray thee, go among us; for it is a stiffnecked people; and pardon our iniquity and our sin,

and take us for thine inheritance. Domine obsecro ut gradiaris nobiscum populus enim durae

ceruicis est et auferas iniquitates nostras atque peccata nos-

34:10 And he said, Behold, I make a covenant: before all que possideas 34:10 respondit Dominus ego inibo pactum

thy people I will do marvels, such as have not been done in all the earth, nor in any nation: and all the people among which thou art shall see the work of the LORD: for it is a terrible thing that I will do with thee.

uidentibus cunctis signa faciam quae numquam sunt uisa super terram nec in ullis gentibus ut cernat populus in cuius

34:11 Observe thou that which I command thee this day: es medio opus Domini terribile quod facturus sum 34:11

behold, I drive out before thee the Amorite, and the Canaanite, and the Hittite, and the Perizzite, and the

Hivite, and the Jebusite. obserua cuncta quae hodie mando tibi ego ipse eiciam ante

faciem tuam Amorreum et Chananeum et Hettheum Fere-

34:12 Take heed to thyself, lest thou make a covenant with zeum quoque et Eueum et Iebuseum 34:12 caue ne umquam the inhabitants of the land whither thou goest, lest it be for a snare in the midst of thee:

cum habitatoribus terrae illius iungas amicitias quae tibi sint

34:13 But ye shall destroy their altars, break their images, in ruinam 34:13 sed aras eorum destrue confringe statuas and cut down their groves:

lucosque succide 34:14 noli adorare deum alienum Domi- 34:14 For thou shalt worship no other god: for the LORD,

whose name is Jealous, is a jealous God:

nus Zelotes nomen eius Deus est aemulator 34:15 ne ineas 34:15 Lest thou make a covenant with the inhabitants of

the land, and they go a whoring after their gods, and do sacrifice unto their gods, and one call thee, and thou eat of

pactum cum hominibus illarum regionum ne cum fornicati his sacrifice; fuerint cum diis suis et adorauerint simulacra eorum uocet

te quispiam ut comedas de immolatis 34:16 nec uxorem de 34:16 And thou take of their daughters unto thy sons, and their daughters go a whoring after their gods, and make thy sons go a whoring after their gods.

filiabus eorum accipies filiis tuis ne postquam ipsae fuerint fornicatae fornicari faciant et filios tuos in deos suos 34:17 34:17 Thou shalt make thee no molten gods. deos conflatiles non facies tibi 34:18 sollemnitatem azymo- 34:18 The feast of unleavened bread shalt thou keep.

Seven days thou shalt eat unleavened bread, as I commanded thee, in the time of the month Abib: for in the

rum custodies septem diebus uesceris azymis sicut praecepi month Abib thou camest out from Egypt. tibi in tempore mensis nouorum mense enim uerni temporis egressus es de Aegypto 34:19 omne quod aperit uuluam 34:19 All that openeth the matrix is mine; and every

firstling among thy cattle, whether ox or sheep, that is male.

generis masculini meum erit de cunctis animantibus tam de bubus quam de ouibus meum erit 34:20 primogenitum asini 34:20 But the firstling of an ass thou shalt redeem with a

lamb: and if thou redeem him not, then shalt thou break his neck. All the firstborn of thy sons thou shalt redeem.

rediies oue sin autem nec pretium pro eo dederis occide- And none shall appear before me empty. tur primogenitum filiorum tuorum rediies nec apparebis in conspectu meo uacuus 34:21 sex diebus operaberis die se- 34:21 Six days thou shalt work, but on the seventh day

thou shalt rest: in earing time and in harvest thou shalt rest.

ptimo cessabis arare et metere 34:22 sollemnitatem ebdo-

34:22 And thou shalt observe the feast of weeks, of the firstfruits of wheat harvest, and the feast of ingathering at the year’s end.

madarum facies tibi in primitiis frugum messis tuae triticeae et sollemnitatem quando redeente anni tempore cuncta con-

duntur 34:23 tribus temporibus anni apparebit omne mas- 34:23 Thrice in the year shall all your menchildren appear

before the LORD God, the God of Israel.

culinum tuum in conspectu omnipotentis Domini Dei Israhel 34:24 cum enim tulero gentes a facie tua et dilatauero termi- 34:24 For I will cast out the nations before thee, and en-

large thy borders: neither shall any man desire thy land, when thou shalt go up to appear before the LORD thy God

nos tuos nullus insidiabitur terrae tuae ascendente te et appa- thrice in the year. rente in conspectu Domini Dei tui ter in anno 34:25 non im- 34:25 Thou shalt not offer the blood of my sacrifice

with leaven; neither shall the sacrifice of the feast of the passover be left unto the morning.

molabis super fermento sanguinem hostiae meae neque residebit mane de uictima sollemnitatis phase 34:26 primitias 34:26 The first of the firstfruits of thy land thou shalt bring

unto the house of the LORD thy God. Thou shalt not seethe a kid in his mother’s milk.

frugum terrae tuae offeres in domum Domini Dei tui non coques hedum in lacte matris suae 34:27 dixitque Dominus ad 34:27 And the LORD said unto Moses, Write thou these

words: for after the tenor of these words I have made a covenant with thee and with Israel.

Mosen scribe tibi uerba haec quibus et tecum et cum Israhel pepigi foedus 34:28 fecit ergo ibi cum Domino quadraginta 34:28 And he was there with the LORD forty days and

forty nights; he did neither eat bread, nor drink water. And he wrote upon the tables the words of the covenant, the ten

dies et quadraginta noctes panem non comedit et aquam non commandments. bibit et scripsit in tabulis uerba foederis decem 34:29 cum- 34:29 And it came to pass, when Moses came down from

mount Sinai with the two tables of testimony in Moses’ hand, when he came down from the mount, that Moses wist not that the skin of his face shone while he talked with him.

que descenderet Moses de monte Sinai tenebat duas tabu-

Biblia Sacra Vulgata

las testimonii et ignorabat quod cornuta esset facies sua ex

34:30 And when Aaron and all the children of Israel saw consortio sermonis Dei 34:30 uidentes autem Aaron et fi-

Moses, behold, the skin of his face shone; and they were afraid to come nigh him.

lii Israhel cornutam Mosi faciem timuerunt prope accedere

34:31 And Moses called unto them; and Aaron and all the 34:31 uocatique ab eo reuersi sunt tam Aaron quam principes

rulers of the congregation returned unto him: and Moses talked with them.

synagogae et postquam locutus est 34:32 uenerunt ad eum

34:32 And afterward all the children of Israel came nigh: and he gave them in commandment all that the LORD had spoken with him in mount Sinai.

etiam omnes filii Israhel quibus praecepit cuncta quae audie-

34:33 And till Moses had done speaking with them, he put rat a Domino in monte Sinai 34:33 impletisque sermonibus

a vail on his face.

34:34 But when Moses went in before the LORD to speak posuit uelamen super faciem suam 34:34 quod ingressus ad

with him, he took the vail off, until he came out. And he came out, and spake unto the children of Israel that which

he was commanded. Dominum et loquens cum eo auferebat donec exiret et tunc

loquebatur ad filios Israhel omnia quae sibi fuerant imperata

34:35 And the children of Israel saw the face of Moses, 34:35 qui uidebant faciem egredientis Mosi esse cornutam

that the skin of Moses’ face shone: and Moses put the vail upon his face again, until he went in to speak with him.

sed operiebat rursus ille faciem suam si quando loquebatur ad eos

35:1 And Moses gathered all the congregation of the chil- 35:1 igitur congregata omni turba filiorum Israhel dixit dren of Israel together, and said unto them, These are the words which the LORD hath commanded, that ye should

do them. ad eos haec sunt quae iussit Dominus fieri 35:2 sex die-

35:2 Six days shall work be done, but on the seventh day there shall be to you an holy day, a sabbath of rest to the LORD: whosoever doeth work therein shall be put to death.

bus facietis opus septimus dies erit uobis sanctus sabbatum

35:3 Ye shall kindle no fire throughout your habitations et requies Domini qui fecerit opus in eo occidetur 35:3 non

upon the sabbath day.

succendetis ignem in omnibus habitaculis uestris per diem

35:4 And Moses spake unto all the congregation of the sabbati 35:4 et ait Moses ad omnem cateruam filiorum Is-

children of Israel, saying, This is the thing which the LORD commanded, saying,

rahel iste est sermo quem praecepit Dominus dicens 35:5

35:5 Take ye from among you an offering unto the LORD: whosoever is of a willing heart, let him bring it, an offering of the LORD; gold, and silver, and brass,

separate apud uos primitias Domino omnis uoluntarius et proni animi offerat eas Domino aurum et argentum et aes

35:6 And blue, and purple, and scarlet, and fine linen, and 35:6 hyacinthum purpuram coccumque bis tinctum et bysgoats’ hair,

35:7 And rams’ skins dyed red, and badgers’ skins, and sum pilos caprarum 35:7 et pelles arietum rubricatas et

shittim wood, 35:8 And oil for the light, and spices for anointing oil, and ianthinas 35:8 ligna setthim 35:9 et oleum ad luminaria for the sweet incense, 35:9 And onyx stones, and stones to be set for the ephod, and for the breastplate.

concinnanda et ut conficiatur unguentum et thymiama sua-

35:10 And every wise hearted among you shall come, and uissimum 35:10 lapides onychinos et gemmas ad ornatum

make all that the LORD hath commanded; 35:11 The tabernacle, his tent, and his covering, his taches, superumeralis et rationalis 35:11 quisquis uestrum est saand his boards, his bars, his pillars, and his sockets,

35:12 The ark, and the staves thereof, with the mercy seat, piens ueniat et faciat quod Dominus imperauit 35:12 taber-

and the vail of the covering,

naculum scilicet et tectum eius atque operimentum anulos

35:13 The table, and his staves, and all his vessels, and the et tabulata cum uectibus paxillos et bases 35:13 arcam et shewbread,

uectes propitiatorium et uelum quod ante illud oppanditur

35:14 The candlestick also for the light, and his furniture, 35:14 mensam cum uectibus et uasis et propositionis panibus and his lamps, with the oil for the light,

35:15 candelabrum ad luminaria sustentanda uasa illius et lu- 35:15 And the incense altar, and his staves, and the anoint-

ing oil, and the sweet incense, and the hanging for the door at the entering in of the tabernacle,

cernas et oleum ad nutrimenta ignium 35:16 altare thymia-

35:16 The altar of burnt offering, with his brasen grate, his staves, and all his vessels, the laver and his foot,

matis et uectes oleum unctionis et thymiama ex aromatibus tentorium ad ostium tabernaculi 35:17 altare holocausti et 35:17 The hangings of the court, his pillars, and their sock-

ets, and the hanging for the door of the court,

craticulam eius aeneam cum uectibus et uasis suis labrum et basim eius 35:18 cortinas atrii cum columnis et basibus 35:18 The pins of the tabernacle, and the pins of the court,

and their cords,

tentorium in foribus uestibuli 35:19 paxillos tabernaculi et 35:19 The cloths of service, to do service in the holy place,

the holy garments for Aaron the priest, and the garments of his sons, to minister in the priest’s office.

atrii cum funiculis suis 35:20 uestimenta quorum usus est

35:20 And all the congregation of the children of Israel departed from the presence of Moses.

in ministerio sanctuarii uestes Aaron pontificis ac filiorum eius ut sacerdotio fungantur mihi 35:21 egressaque omnis 35:21 And they came, every one whose heart stirred him

up, and every one whom his spirit made willing, and they brought the LORD’s offering to the work of the tabernacle of the congregation, and for all his service, and for the holy garments.

multitudo filiorum Israhel de conspectu Mosi 35:22 obtu-

35:22 And they came, both men and women, as many as were willing hearted, and brought bracelets, and earrings, and rings, and tablets, all jewels of gold: and every man that offered offered an offering of gold unto the LORD.

lit mente promptissima atque deuota primitias Domino ad faciendum opus tabernaculi testimonii quicquid in cultum et ad uestes sanctas necessarium erat 35:23 uiri cum mu- 35:23 And every man, with whom was found blue, and

purple, and scarlet, and fine linen, and goats’ hair, and red skins of rams, and badgers’ skins, brought them.

lieribus praebuerunt armillas et inaures anulos et dextralia omne uas aureum in donaria Domini separatum est 35:24 35:24 Every one that did offer an offering of silver and

brass brought the LORD’s offering: and every man, with whom was found shittim wood for any work of the service,

si quis habuit hyacinthum purpuram coccumque bis tinctum brought it. byssum et pilos caprarum pelles arietum rubricatas et ianthinas 35:25 argenti et aeris metalla obtulerunt Domino li- 35:25 And all the women that were wise hearted did spin

with their hands, and brought that which they had spun, both of blue, and of purple, and of scarlet, and of fine

gnaque setthim in uarios usus 35:26 sed et mulieres doc- linen.

35:26 And all the women whose heart stirred them up in wisdom spun goats’ hair.

tae dederunt quae neuerant hyacinthum purpuram et uermiculum ac byssum 35:27 et pilos caprarum sponte propria 35:27 And the rulers brought onyx stones, and stones to

be set, for the ephod, and for the breastplate;

cuncta tribuentes 35:28 principes uero obtulerunt lapides 35:28 And spice, and oil for the light, and for the anointing

oil, and for the sweet incense.

onychinos et gemmas ad superumerale et rationale 35:29 35:29 The children of Israel brought a willing offering

unto the LORD, every man and woman, whose heart made them willing to bring for all manner of work, which the

aromataque et oleum ad luminaria concinnanda et ad prae- LORD had commanded to be made by the hand of Moses. parandum unguentum ac thymiama odoris suauissimi con-

ponendum 35:30 omnes uiri et mulieres mente deuota ob- 35:30 And Moses said unto the children of Israel, See, the LORD hath called by name Bezaleel the son of Uri, the son of Hur, of the tribe of Judah;

tulerunt donaria ut fierent opera quae iusserat Dominus per manum Mosi cuncti filii Israhel uoluntaria Domino dedicauerunt 35:31 dixitque Moses ad filios Israhel ecce uocauit 35:31 And he hath filled him with the spirit of God, in

wisdom, in understanding, and in knowledge, and in all manner of workmanship;

Dominus ex nomine Beselehel filium Vri filii Hur de tribu Iuda 35:32 impleuitque eum spiritu Dei sapientiae et intel- 35:32 And to devise curious works, to work in gold, and

in silver, and in brass,

legentiae et scientiae omni doctrina 35:33 ad excogitandum 35:33 And in the cutting of stones, to set them, and in

carving of wood, to make any manner of cunning work.

Biblia Sacra Vulgata

et faciendum opus in auro et argento et aere sculpendisque lapidibus et opere carpentario quicquid fabre adinueniri po-

35:34 And he hath put in his heart that he may teach, both test 35:34 dedit in corde eius Hooliab quoque filium Achi-

he, and Aholiab, the son of Ahisamach, of the tribe of Dan.

35:35 Them hath he filled with wisdom of heart, to work samech de tribu Dan 35:35 ambos erudiuit sapientia ut fa-

all manner of work, of the engraver, and of the cunning workman, and of the embroiderer, in blue, and in purple, in scarlet, and in fine linen, and of the weaver, even of them that do any work, and of those that devise cunning work.

ciant opera abietarii polymitarii ac plumarii de hyacintho et purpura coccoque bis tincto et bysso et texant omnia ac noua quaeque repperiant

36:1 Then wrought Bezaleel and Aholiab, and every wise 36:1 fecit ergo Beselehel et Hooliab et omnis uir sapiens

hearted man, in whom the LORD put wisdom and understanding to know how to work all manner of work for the service of the sanctuary, according to all that the LORD had commanded.

quibus dedit Dominus sapientiam et intellectum ut scirent fabre operari quae in usus sanctuarii necessaria sunt et quae

36:2 And Moses called Bezaleel and Aholiab, and every praecepit Dominus 36:2 cumque uocasset eos Moses et wise hearted man, in whose heart the LORD had put wisdom, even every one whose heart stirred him up to come

unto the work to do it: omnem eruditum uirum cui dederat Deus sapientiam et qui 36:3 And they received of Moses all the offering, which sponte sua obtulerant se ad faciendum opus 36:3 tradidit

the children of Israel had brought for the work of the service of the sanctuary, to make it withal. And they brought

yet unto him free offerings every morning. eis uniuersa donaria filiorum Israhel qui cum instarent operi 36:4 And all the wise men, that wrought all the work of cotidie mane uota populus offerebat 36:4 unde artifices ue-

the sanctuary, came every man from his work which they made;

nire conpulsi 36:5 dixerunt Mosi plus offert populus quam

36:5 And they spake unto Moses, saying, The people bring much more than enough for the service of the work, which the LORD commanded to make.

necessarium est 36:6 iussit ergo Moses praeconis uoce can-

36:6 And Moses gave commandment, and they caused it to be proclaimed throughout the camp, saying, Let neither man nor woman make any more work for the offering of the sanctuary. So the people were restrained from bringing.

tari nec uir nec mulier quicquam ultra offerat in opere sanctuarii sicque cessatum est a muneribus offerendis 36:7 eo

36:7 For the stuff they had was sufficient for all the work to make it, and too much.

36:8 And every wise hearted man among them that quod oblata sufficerent et superabundarent 36:8 fecerunt-

wrought the work of the tabernacle made ten curtains of fine twined linen, and blue, and purple, and scarlet: with

cherubims of cunning work made he them. que omnes corde sapientes ad explendum opus tabernaculi

cortinas decem de bysso retorta et hyacintho et purpura coc-

36:9 The length of one curtain was twenty and eight cu- coque bis tincto opere uario et arte polymita 36:9 quarum

bits, and the breadth of one curtain four cubits: the curtains were all of one size.

una habebat in longitudine uiginti octo cubitos et in latitu-

36:10 And he coupled the five curtains one unto another: dine quattuor una mensura erat omnium cortinarum 36:10

and the other five curtains he coupled one unto another.

coniunxitque cortinas quinque alteram alteri et alias quinque

36:11 And he made loops of blue on the edge of one cur- sibi inuicem copulauit 36:11 fecit et ansas hyacinthinas in

tain from the selvedge in the coupling: likewise he made in the uttermost side of another curtain, in the coupling of

the second. ora cortinae unius ex utroque latere et in ora cortinae alte-36:12 Fifty loops made he in one curtain, and fifty loops rius similiter 36:12 ut contra se inuicem uenirent ansae et

made he in the edge of the curtain which was in the coupling of the second: the loops held one curtain to another.

mutuo iungerentur 36:13 unde et quinquaginta fudit circu-

36:13 And he made fifty taches of gold, and coupled the curtains one unto another with the taches: so it became one tabernacle.

los aureos qui morderent cortinarum ansas et fieret unum ta-

36:14 And he made curtains of goats’ hair for the tent over bernaculum 36:14 fecit et saga undecim de pilis caprarum the tabernacle: eleven curtains he made them. 36:15 The length of one curtain was thirty cubits, and four ad operiendum tectum tabernaculi 36:15 unum sagum hacubits was the breadth of one curtain: the eleven curtains were of one size.

bebat in longitudine cubitos triginta et in latitudine cubitos

quattuor unius mensurae erant omnia saga 36:16 quorum 36:16 And he coupled five curtains by themselves, and six

curtains by themselves.

quinque iunxit seorsum et sex alia separatim 36:17 fecit- 36:17 And he made fifty loops upon the uttermost edge of

the curtain in the coupling, and fifty loops made he upon the edge of the curtain which coupleth the second.

que ansas quinquaginta in ora sagi unius et quinquaginta in ora sagi alterius ut sibi inuicem iungerentur 36:18 et fibu- 36:18 And he made fifty taches of brass to couple the tent

together, that it might be one.

las aeneas quinquaginta quibus necteretur tectum et unum pallium ex omnibus sagis fieret 36:19 fecit et opertorium 36:19 And he made a covering for the tent of rams’ skins

dyed red, and a covering of badgers’ skins above that.

tabernaculi de pellibus arietum rubricatis aliudque desuper uelamentum de pellibus ianthinis 36:20 fecit et tabulas ta- 36:20 And he made boards for the tabernacle of shittim

wood, standing up.

bernaculi de lignis setthim stantes 36:21 decem cubitorum 36:21 The length of a board was ten cubits, and the breadth

of a board one cubit and a half.

erat longitudo tabulae unius et unum ac semis cubitum latitudo retinebat 36:22 binae incastraturae erant per singulas 36:22 One board had two tenons, equally distant one from

another: thus did he make for all the boards of the tabernacle.

tabulas ut altera alteri iungeretur sic fecit in omnibus tabulis tabernaculi 36:23 e quibus uiginti ad plagam meridia- 36:23 And he made boards for the tabernacle; twenty

boards for the south side southward:

nam erant contra austrum 36:24 cum quadraginta basibus 36:24 And forty sockets of silver he made under the

twenty boards; two sockets under one board for his two tenons, and two sockets under another board for his two

argenteis duae bases sub una tabula ponebantur ex utraque tenons. angulorum parte ubi incastraturae laterum in angulis terminantur 36:25 ad plagam quoque tabernaculi quae respicit ad 36:25 And for the other side of the tabernacle, which is

toward the north corner, he made twenty boards,

aquilonem fecit uiginti tabulas 36:26 cum quadraginta ar- 36:26 And their forty sockets of silver; two sockets under

one board, and two sockets under another board.

genteis basibus duas bases per singulas tabulas 36:27 con- 36:27 And for the sides of the tabernacle westward he

made six boards.

tra occidentem uero id est ad eam partem tabernaculi quae mare respicit fecit sex tabulas 36:28 et duas alias per singu- 36:28 And two boards made he for the corners of the taber-

nacle in the two sides.

los angulos tabernaculi retro 36:29 quae iunctae erant deor- 36:29 And they were coupled beneath, and coupled to-

gether at the head thereof, to one ring: thus he did to both of them in both the corners.

sum usque sursum et in unam conpagem pariter ferebantur ita fecit ex utraque parte per angulos 36:30 ut octo essent 36:30 And there were eight boards; and their sockets were

sixteen sockets of silver, under every board two sockets.

simul tabulae et haberent bases argenteas sedecim binas scilicet bases sub singulis tabulis 36:31 fecit et uectes de lignis 36:31 And he made bars of shittim wood; five for the

boards of the one side of the tabernacle,

setthim quinque ad continendas tabulas unius lateris tabernaculi 36:32 et quinque alios ad alterius lateris tabulas co- 36:32 And five bars for the boards of the other side of the

tabernacle, and five bars for the boards of the tabernacle for the sides westward.

aptandas et extra hos quinque alios uectes ad occidentalem plagam tabernaculi contra mare 36:33 fecit quoque uectem 36:33 And he made the middle bar to shoot through the

boards from the one end to the other.

alium qui per medias tabulas ab angulo usque ad angulum perueniret 36:34 ipsa autem tabulata deaurauit et circulos 36:34 And he overlaid the boards with gold, and made

their rings of gold to be places for the bars, and overlaid the bars with gold.

Biblia Sacra Vulgata

eorum fecit aureos per quos uectes induci possint quos et

36:35 And he made a vail of blue, and purple, and scar- ipsos aureis lamminis operuit 36:35 fecit et uelum de hya-

let, and fine twined linen: with cherubims made he it of cunning work.

cintho purpura uermiculo ac bysso retorta opere polymitario

36:36 And he made thereunto four pillars of shittim wood, uarium atque distinctum 36:36 et quattuor columnas de lig-

and overlaid them with gold: their hooks were of gold; and he cast for them four sockets of silver.

nis setthim quas cum capitibus deaurauit fusis basibus earum

36:37 And he made an hanging for the tabernacle door argenteis 36:37 fecit et tentorium in introitu tabernaculi ex

of blue, and purple, and scarlet, and fine twined linen, of needlework;

hyacintho purpura uermiculo byssoque retorta opere pluma-

36:38 And the five pillars of it with their hooks: and he rii 36:38 et columnas quinque cum capitibus suis quas ope-

overlaid their chapiters and their fillets with gold: but their five sockets were of brass.

ruit auro basesque earum fudit aeneas

37:1 And Bezaleel made the ark of shittim wood: two cu- 37:1 fecit autem Beselehel et arcam de lignis setthim ha-

bits and a half was the length of it, and a cubit and a half the breadth of it, and a cubit and a half the height of it:

bentem duos semis cubitos in longitudinem et cubitum ac semissem in latitudinem altitudo quoque uno cubito fuit et

37:2 And he overlaid it with pure gold within and without, dimidio uestiuitque eam auro purissimo intus ac foris 37:2

and made a crown of gold to it round about.

37:3 And he cast for it four rings of gold, to be set by the et fecit illi coronam auream per gyrum 37:3 conflans quat-

four corners of it; even two rings upon the one side of it, and two rings upon the other side of it.

tuor anulos aureos per quattuor angulos eius duos anulos in

37:4 And he made staves of shittim wood, and overlaid latere uno et duos in altero 37:4 uectes quoque fecit de li-

them with gold.

37:5 And he put the staves into the rings by the sides of gnis setthim quos uestiuit auro 37:5 et quos misit in anu-

the ark, to bear the ark.

37:6 And he made the mercy seat of pure gold: two cubits los qui erant in lateribus arcae ad portandum eam 37:6 fe-

and a half was the length thereof, and one cubit and a half the breadth thereof.

cit et propitiatorium id est oraculum de auro mundissimo duorum cubitorum et dimidio in longitudine et cubito ac se-

37:7 And he made two cherubims of gold, beaten out of misse in latitudine 37:7 duos etiam cherubin ex auro duc-

one piece made he them, on the two ends of the mercy seat;

tili quos posuit ex utraque parte propitiatorii 37:8 cherub

37:8 One cherub on the end on this side, and another cherub on the other end on that side: out of the mercy seat made he the cherubims on the two ends thereof.

unum in summitate huius partis et cherub alterum in summitate partis alterius duos cherubin in singulis summitati-

37:9 And the cherubims spread out their wings on high, bus propitiatorii 37:9 extendentes alas et tegentes propitia-

and covered with their wings over the mercy seat, with their faces one to another; even to the mercy seatward were

the faces of the cherubims. torium seque mutuo et illud respectantes 37:10 fecit et men-

37:10 And he made the table of shittim wood: two cubits was the length thereof, and a cubit the breadth thereof, and a cubit and a half the height thereof:

sam de lignis setthim in longitudine duorum cubitorum et in latitudine unius cubiti quae habebat in altitudine cubitum ac

37:11 And he overlaid it with pure gold, and made there- semissem 37:11 circumdeditque eam auro mundissimo et unto a crown of gold round about.

37:12 Also he made thereunto a border of an handbreadth fecit illi labium aureum per gyrum 37:12 ipsique labio co-

round about; and made a crown of gold for the border thereof round about.

ronam interrasilem quattuor digitorum et super eandem al-

37:13 And he cast for it four rings of gold, and put the teram coronam auream 37:13 fudit et quattuor circulos au-

rings upon the four corners that were in the four feet thereof.

reos quos posuit in quattuor angulis per singulos pedes mensae 37:14 contra coronam misitque in eos uectes ut possit 37:14 Over against the border were the rings, the places

for the staves to bear the table.

mensa portari 37:15 ipsos quoque uectes fecit de lignis set- 37:15 And he made the staves of shittim wood, and over-

laid them with gold, to bear the table.

thim et circumdedit eos auro 37:16 et uasa ad diuersos usus 37:16 And he made the vessels which were upon the table,

his dishes, and his spoons, and his bowls, and his covers to cover withal, of pure gold.

mensae acetabula fialas cyatos et turibula ex auro puro in quibus offerenda sunt liba 37:17 fecit et candelabrum duc- 37:17 And he made the candlestick of pure gold: of beaten

work made he the candlestick; his shaft, and his branch, his bowls, his knops, and his flowers, were of the same:

tile de auro mundissimo de cuius uecte calami scyphi spherulae ac lilia procedebant 37:18 sex in utroque latere tres 37:18 And six branches going out of the sides thereof;

three branches of the candlestick out of the one side thereof, and three branches of the candlestick out of the

calami ex parte una et tres ex altera 37:19 tres scyphi in other side thereof:

37:19 Three bowls made after the fashion of almonds in one branch, a knop and a flower; and three bowls made like almonds in another branch, a knop and a flower: so throughout the six branches going out of the candlestick.

nucis modum per calamos singulos spherulaeque simul et lilia et tres scyphi instar nucis in calamo altero spherulaeque simul et lilia aequum erat opus sex calamorum qui procedebant de stipite candelabri 37:20 in ipso autem uecte erant 37:20 And in the candlestick were four bowls made like

almonds, his knops, and his flowers:

quattuor scyphi in nucis modum spherulaeque per singulos et lilia 37:21 et spherae sub duobus calamis per loca tria 37:21 And a knop under two branches of the same, and a

knop under two branches of the same, and a knop under two branches of the same, according to the six branches

qui simul sex fiunt calami procedentes de uecte uno 37:22 going out of it.

37:22 Their knops and their branches were of the same: all of it was one beaten work of pure gold.

et spherae igitur et calami ex ipso erant uniuersa ductilia de auro purissimo 37:23 fecit et lucernas septem cum emunc- 37:23 And he made his seven lamps, and his snuffers, and

his snuffdishes, of pure gold.

toriis suis et uasa ubi quae emuncta sunt extinguuntur de auro mundissimo 37:24 talentum auri adpendebat candela- 37:24 Of a talent of pure gold made he it, and all the ves-

sels thereof.

brum cum omnibus uasis suis 37:25 fecit et altare thymia- 37:25 And he made the incense altar of shittim wood: the

length of it was a cubit, and the breadth of it a cubit; it was foursquare; and two cubits was the height of it; the horns

matis de lignis setthim habens per quadrum singulos cubi- thereof were of the same. tos et in altitudine duos e cuius angulis procedebant cornua 37:26 uestiuitque illud auro purissimo cum craticula ac parie- 37:26 And he overlaid it with pure gold, both the top of it,

and the sides thereof round about, and the horns of it: also he made unto it a crown of gold round about.

tibus et cornibus 37:27 fecitque ei coronam aureolam per

37:27 And he made two rings of gold for it under the crown thereof, by the two corners of it, upon the two sides thereof, to be places for the staves to bear it withal.

gyrum et duos anulos aureos sub corona per singula latera ut mittantur in eos uectes et possit altare portari 37:28 ip- 37:28 And he made the staves of shittim wood, and over-

laid them with gold.

sos autem uectes fecit de lignis setthim et operuit lamminis aureis 37:29 conposuit et oleum ad sanctificationis unguen- 37:29 And he made the holy anointing oil, and the pure in-

cense of sweet spices, according to the work of the apothecary.

tum et thymiama de aromatibus mundissimis opere pigmentarii

38:1 fecit et altare holocausti de lignis setthim quinque cu- 38:1 And he made the altar of burnt offering of shittim

wood: five cubits was the length thereof, and five cubits the breadth thereof; it was foursquare; and three cubits the

bitorum per quadrum et trium in altitudine 38:2 cuius cor- height thereof.

38:2 And he made the horns thereof on the four corners of it; the horns thereof were of the same: and he overlaid it with brass.

nua de angulis procedebant operuitque illud aeneis lammi-

Biblia Sacra Vulgata

38:3 And he made all the vessels of the altar, the pots, and nis 38:3 et in usus eius parauit ex aere uasa diuersa lebethe shovels, and the basons, and the fleshhooks, and the firepans: all the vessels thereof made he of brass.

tas forcipes fuscinulas uncinos et ignium receptacula 38:4

38:4 And he made for the altar a brasen grate of network under the compass thereof beneath unto the midst of it.

craticulamque eius in modum retis fecit aeneam et subter

38:5 And he cast four rings for the four ends of the grate eam in altaris medio arulam 38:5 fusis quattuor anulis per

of brass, to be places for the staves.

totidem retiaculi summitates ad inmittendos uectes ad por-

38:6 And he made the staves of shittim wood, and overlaid tandum 38:6 quos et ipsos fecit de lignis setthim et operuit

them with brass.

38:7 And he put the staves into the rings on the sides of lamminis aeneis 38:7 induxitque in circulos qui in altaris

the altar, to bear it withal; he made the altar hollow with boards.

lateribus eminebant ipsum autem altare non erat solidum sed

38:8 And he made the laver of brass, and the foot of it cauum ex tabulis et intus uacuum 38:8 fecit et labrum ae-

of brass, of the lookingglasses of the women assembling, which assembled at the door of the tabernacle of the con-

gregation. neum cum base sua de speculis mulierum quae excubabant 38:9 And he made the court: on the south side southward in ostio tabernaculi 38:9 et atrium in cuius australi plaga

the hangings of the court were of fine twined linen, an hundred cubits:

erant tentoria de bysso retorta cubitorum centum 38:10 co-

38:10 Their pillars were twenty, and their brasen sockets twenty; the hooks of the pillars and their fillets were of silver.

lumnae aeneae uiginti cum basibus suis capita columnarum

38:11 And for the north side the hangings were an hundred et tota operis celatura argentea 38:11 aeque ad septentrio-

cubits, their pillars were twenty, and their sockets of brass twenty; the hooks of the pillars and their fillets of silver.

nalis plagam tentoria columnae basesque et capita columna-

38:12 And for the west side were hangings of fifty cubits, rum eiusdem et mensurae et operis ac metalli erant 38:12 their pillars ten, and their sockets ten; the hooks of the pillars and their fillets of silver.

in ea uero plaga quae occidentem respicit fuere tentoria cubitorum quinquaginta columnae decem cum basibus suis ae-

38:13 And for the east side eastward fifty cubits. neae et capita columnarum celata argentea 38:13 porro con-38:14 The hangings of the one side of the gate were fifteen tra orientem quinquaginta cubitorum parauit tentoria 38:14

cubits; their pillars three, and their sockets three.

e quibus quindecim cubitos columnarum trium cum basibus

38:15 And for the other side of the court gate, on this hand suis unum tenebat latus 38:15 et in parte altera quia utra-

and that hand, were hangings of fifteen cubits; their pillars three, and their sockets three.

que introitum tabernaculi facit quindecim aeque cubitorum

38:16 All the hangings of the court round about were of erant tentoria columnae tres et bases totidem 38:16 cuncta

fine twined linen.

38:17 And the sockets for the pillars were of brass; the atrii tentoria byssus torta texuerat 38:17 bases columnarum

hooks of the pillars and their fillets of silver; and the overlaying of their chapiters of silver; and all the pillars of the

court were filleted with silver. fuere aeneae capita autem earum cum celaturis suis argentea 38:18 And the hanging for the gate of the court was sed et ipsas columnas atrii uestiuit argento 38:18 et in intro-

needlework, of blue, and purple, and scarlet, and fine twined linen: and twenty cubits was the length, and the height in the breadth was five cubits, answerable to the hangings of the court.

itu eius opere plumario fecit tentorium ex hyacintho purpura uermiculo ac bysso retorta quod habebat uiginti cubitos in longitudine altitudo uero quinque cubitorum erat iuxta men-

38:19 And their pillars were four, and their sockets of suram quam cuncta atrii habebant tentoria 38:19 columnae

brass four; their hooks of silver, and the overlaying of their chapiters and their fillets of silver.

autem ingressus fuere quattuor cum basibus aeneis capita-

38:20 And all the pins of the tabernacle, and of the court que earum et celaturae argenteae 38:20 paxillos quoque ta-

round about, were of brass.

bernaculi et atrii per gyrum fecit aeneos 38:21 haec sunt 38:21 This is the sum of the tabernacle, even of the taber-

nacle of testimony, as it was counted, according to the commandment of Moses, for the service of the Levites,

instrumenta tabernaculi testimonii quae numerata sunt iuxta by the hand of Ithamar, son to Aaron the priest. praeceptum Mosi in caerimonias Leuitarum per manum Ithamar filii Aaron sacerdotis 38:22 quas Beselehel filius Vri 38:22 And Bezaleel the son Uri, the son of Hur, of the tribe

of Judah, made all that the LORD commanded Moses.

filii Hur de tribu Iuda Domino per Mosen iubente conpleuerat 38:23 iuncto sibi socio Hooliab filio Achisamech de 38:23 And with him was Aholiab, son of Ahisamach, of

the tribe of Dan, an engraver, and a cunning workman, and an embroiderer in blue, and in purple, and in scarlet,

tribu Dan qui et ipse artifex lignorum egregius fuit et po- and fine linen. lymitarius atque plumarius ex hyacintho purpura uermiculo

et bysso 38:24 omne aurum quod expensum est in opere 38:24 All the gold that was occupied for the work in all the work of the holy place, even the gold of the offering, was twenty and nine talents, and seven hundred and thirty

sanctuarii et quod oblatum in donariis uiginti nouem talen- shekels, after the shekel of the sanctuary. torum fuit et septingentorum triginta siclorum ad mensuram sanctuarii 38:25 oblatum est autem ab his qui transierant 38:25 And the silver of them that were numbered of

the congregation was an hundred talents, and a thousand seven hundred and threescore and fifteen shekels, after the

ad numerum a uiginti annis et supra de sescentis tribus mi- shekel of the sanctuary: libus et quingentis quinquaginta armatorum 38:26 fuerunt 38:26 A bekah for every man, that is, half a shekel, af-

ter the shekel of the sanctuary, for every one that went to be numbered, from twenty years old and upward, for six hundred thousand and three thousand and five hundred and fifty men.

praeterea centum talenta argenti e quibus conflatae sunt bases sanctuarii et introitus ubi uelum pendet 38:27 centum 38:27 And of the hundred talents of silver were cast the

sockets of the sanctuary, and the sockets of the vail; an hundred sockets of the hundred talents, a talent for a

bases factae sunt de talentis centum singulis talentis per ba- socket. ses singulas supputatis 38:28 de mille autem septingentis et 38:28 And of the thousand seven hundred seventy and five

shekels he made hooks for the pillars, and overlaid their chapiters, and filleted them.

septuaginta quinque fecit capita columnarum quas et ipsas uestiuit argento 38:29 aeris quoque oblata sunt talenta sep- 38:29 And the brass of the offering was seventy talents,

and two thousand and four hundred shekels.

tuaginta duo milia et quadringenti supra sicli 38:30 ex qui- 38:30 And therewith he made the sockets to the door of

the tabernacle of the congregation, and the brasen altar, and the brasen grate for it, and all the vessels of the altar,

bus fusae sunt bases in introitu tabernaculi testimonii et altare aeneum cum craticula sua omniaque uasa quae ad usum eius pertinent 38:31 et bases atrii tam in circuitu quam in 38:31 And the sockets of the court round about, and the

sockets of the court gate, and all the pins of the tabernacle, and all the pins of the court round about.

ingressu eius et paxilli tabernaculi atque atrii per gyrum

39:1 de hyacintho uero et purpura uermiculo ac bysso fecit 39:1 And of the blue, and purple, and scarlet, they made

cloths of service, to do service in the holy place, and made the holy garments for Aaron; as the LORD commanded

uestes quibus indueretur Aaron quando ministrabat in sanc- Moses. tis sicut praecepit Dominus Mosi 39:2 fecit igitur superu- 39:2 And he made the ephod of gold, blue, and purple, and

scarlet, and fine twined linen.

merale de auro hyacintho et purpura coccoque bis tincto et bysso retorta 39:3 opere polymitario inciditque bratteas au- 39:3 And they did beat the gold into thin plates, and cut it

into wires, to work it in the blue, and in the purple, and in the scarlet, and in the fine linen, with cunning work.

reas et extenuauit in fila ut possint torqueri cum priorum colorum subtemine 39:4 duasque oras sibi inuicem copulatas 39:4 They made shoulderpieces for it, to couple it to-

gether: by the two edges was it coupled together.

in utroque latere summitatum 39:5 et balteum ex hisdem 39:5 And the curious girdle of his ephod, that was upon it, was of the same, according to the work thereof; of gold, blue, and purple, and scarlet, and fine twined linen; as the LORD commanded Moses.

Biblia Sacra Vulgata

39:6 And they wrought onyx stones inclosed in ouches of coloribus sicut praeceperat Dominus Mosi 39:6 parauit et

gold, graven, as signets are graven, with the names of the children of Israel.

duos lapides onychinos adstrictos et inclusos auro et sculptos

39:7 And he put them on the shoulders of the ephod, that arte gemmaria nominibus filiorum Israhel 39:7 posuitque

they should be stones for a memorial to the children of Israel; as the LORD commanded Moses.

eos in lateribus superumeralis in monumentum filiorum Isra-

39:8 And he made the breastplate of cunning work, like hel sicut praeceperat Dominus Mosi 39:8 fecit et rationale

the work of the ephod; of gold, blue, and purple, and scarlet, and fine twined linen.

opere polymito iuxta opus superumeralis ex auro hyacintho

39:9 It was foursquare; they made the breastplate dou- purpura coccoque bis tincto et bysso retorta 39:9 quadran-

ble: a span was the length thereof, and a span the breadth thereof, being doubled.

gulum duplex mensurae palmi 39:10 et posuit in eo gem-

39:10 And they set in it four rows of stones: the first row was a sardius, a topaz, and a carbuncle: this was the first row.

marum ordines quattuor in primo uersu erat sardius topazius

39:11 And the second row, an emerald, a sapphire, and a zmaragdus 39:11 in secundo carbunculus sapphyrus iaspis

diamond.

39:12 And the third row, a ligure, an agate, and an 39:12 in tertio ligyrius achates amethistus 39:13 in quarto

amethyst. 39:13 And the fourth row, a beryl, an onyx, and a jasper: they were inclosed in ouches of gold in their inclosings.

chrysolitus onychinus berillus circumdati et inclusi auro per

39:14 And the stones were according to the names of the ordines suos 39:14 ipsique lapides duodecim sculpti erant

children of Israel, twelve, according to their names, like the engravings of a signet, every one with his name, ac-

cording to the twelve tribes. nominibus duodecim tribuum Israhel singuli per nomina sin-39:15 And they made upon the breastplate chains at the gulorum 39:15 fecerunt in rationali et catenulas sibi inui-

ends, of wreathen work of pure gold.

39:16 And they made two ouches of gold, and two gold cem coherentes de auro purissimo 39:16 et duos uncinos

rings; and put the two rings in the two ends of the breastplate.

totidemque anulos aureos porro anulos posuerunt in utro-

39:17 And they put the two wreathen chains of gold in the que latere rationalis 39:17 e quibus penderent duae catenae

two rings on the ends of the breastplate.

aureae quas inseruerunt uncinis qui in superumeralis angu-

39:18 And the two ends of the two wreathen chains they lis eminebant 39:18 haec et ante et retro ita conueniebant

fastened in the two ouches, and put them on the shoulderpieces of the ephod, before it.

sibi ut superumerale et rationale mutuo necterentur 39:19

39:19 And they made two rings of gold, and put them on the two ends of the breastplate, upon the border of it, which was on the side of the ephod inward.

stricta ad balteum et anulis fortius copulata quos iungebat uitta hyacinthina ne laxe fluerent et a se inuicem mouerentur

39:20 And they made two other golden rings, and put sicut praecepit Dominus Mosi 39:20 fecerunt quoque tuni-

them on the two sides of the ephod underneath, toward the forepart of it, over against the other coupling thereof,

above the curious girdle of the ephod. cam superumeralis totam hyacinthinam 39:21 et capitium

39:21 And they did bind the breastplate by his rings unto the rings of the ephod with a lace of blue, that it might be above the curious girdle of the ephod, and that the breastplate might not be loosed from the ephod; as the LORD commanded Moses.

in superiori parte contra medium oramque per gyrum capitii textilem 39:22 deorsum autem ad pedes mala punica ex

39:22 And he made the robe of the ephod of woven work, all of blue.

hyacintho purpura uermiculo ac bysso retorta 39:23 et tin-

39:23 And there was an hole in the midst of the robe, as the hole of an habergeon, with a band round about the hole, that it should not rend.

tinabula de auro mundissimo quae posuerunt inter mala gra-

39:24 And they made upon the hems of the robe nata in extrema parte tunicae per gyrum 39:24 tintinabulum

pomegranates of blue, and purple, and scarlet, and twined linen.

aureum et malum punicum quibus ornatus incedebat pontifex quando ministerio fungebatur sicut praecepit Dominus

39:25 And they made bells of pure gold, and put the Mosi 39:25 fecerunt et tunicas byssinas opere textili Aaron

bells between the pomegranates upon the hem of the robe, round about between the pomegranates;

et filiis eius 39:26 et mitras cum coronulis suis ex bysso 39:26 A bell and a pomegranate, a bell and a pomegranate,

round about the hem of the robe to minister in; as the LORD commanded Moses.

39:27 feminalia quoque linea byssina 39:28 cingulum uero

39:27 And they made coats of fine linen of woven work for Aaron, and for his sons, 39:28 And a mitre of fine linen, and goodly bonnets of fine linen, and linen breeches of fine twined linen,

de bysso retorta hyacintho purpura ac uermiculo distinctum arte plumaria sicut praecepit Dominus Mosi 39:29 fecerunt 39:29 And a girdle of fine twined linen, and blue, and pur-

ple, and scarlet, of needlework; as the LORD commanded Moses.

et lamminam sacrae uenerationis de auro purissimo scripseruntque in ea opere gemmario Sanctum Domini 39:30 et 39:30 And they made the plate of the holy crown of pure

gold, and wrote upon it a writing, like to the engravings of a signet, HOLINESS TO THE LORD.

strinxerunt eam cum mitra uitta hyacinthina sicut praecepit Dominus Mosi 39:31 perfectum est igitur omne opus taber- 39:31 And they tied unto it a lace of blue, to fasten it on

high upon the mitre; as the LORD commanded Moses.

naculi et tecti testimonii feceruntque filii Israhel cuncta quae praeceperat Dominus Mosi 39:32 et obtulerunt tabernacu- 39:32 Thus was all the work of the tabernacle of the tent

of the congregation finished: and the children of Israel did according to all that the LORD commanded Moses, so did

lum et tectum et uniuersam supellectilem anulos tabulas uec- they. tes columnas ac bases 39:33 opertorium de pellibus arietum 39:33 And they brought the tabernacle unto Moses, the

tent, and all his furniture, his taches, his boards, his bars, and his pillars, and his sockets,

rubricatis et aliud operimentum de ianthinis pellibus 39:34

39:34 And the covering of rams’ skins dyed red, and the covering of badgers’ skins, and the vail of the covering,

uelum arcam uectes propitiatorium 39:35 mensam cum ua- 39:35 The ark of the testimony, and the staves thereof, and

the mercy seat,

sis et propositionis panibus 39:36 candelabrum lucernas et 39:36 The table, and all the vessels thereof, and the shew-

bread,

utensilia eorum cum oleo 39:37 altare aureum et unguen- 39:37 The pure candlestick, with the lamps thereof, even with the lamps to be set in order, and all the vessels thereof, and the oil for light,

tum thymiama ex aromatibus 39:38 et tentorium in introitu

39:38 And the golden altar, and the anointing oil, and the sweet incense, and the hanging for the tabernacle door,

tabernaculi 39:39 altare aeneum retiaculum uectes et uasa _{39:39 The brasen altar, and his grate of brass, his staves,}

and all his vessels, the laver and his foot,

eius omnia labrum cum basi sua tentoria atrii et columnas cum basibus suis 39:40 tentorium in introitu atrii funicu- 39:40 The hangings of the court, his pillars, and his sock-

ets, and the hanging for the court gate, his cords, and his pins, and all the vessels of the service of the tabernacle,

losque illius et paxillos nihil ex uasis defuit quae in ministe- for the tent of the congregation, rium tabernaculi et in tectum foederis iussa sunt fieri 39:41 39:41 The cloths of service to do service in the holy place,

and the holy garments for Aaron the priest, and his sons’ garments, to minister in the priest’s office.

uestes quoque quibus sacerdotes utuntur in sanctuario Aaron scilicet et filii eius 39:42 obtulerunt filii Israhel sicut prae- 39:42 According to all that the LORD commanded Moses,

so the children of Israel made all the work.

ceperat Dominus 39:43 quae postquam Moses cuncta uidit 39:43 And Moses did look upon all the work, and, behold,

they had done it as the LORD had commanded, even so had they done it: and Moses blessed them.

expleta benedixit eis

40:1 locutusque est Dominus ad Mosen dicens 40:2 40:1 And the LORD spake unto Moses, saying,

40:2 On the first day of the first month shalt thou set up the tabernacle of the tent of the congregation.

mense primo die prima mensis eriges tabernaculum testimo-

nii 40:3 et pones in eo arcam dimittesque ante illam ue- 40:3 And thou shalt put therein the ark of the testimony,

and cover the ark with the vail.

lum 40:4 et inlata mensa pones super eam quae rite prae- 40:4 And thou shalt bring in the table, and set in order the things that are to be set in order upon it; and thou shalt bring in the candlestick, and light the lamps thereof.

cepta sunt candelabrum stabit cum lucernis suis 40:5 et al-

40:5 And thou shalt set the altar of gold for the incense before the ark of the testimony, and put the hanging of the door to the tabernacle.

tare aureum in quo adoletur incensum coram arca testimonii tentorium in introitu tabernaculi pones 40:6 et ante illud 40:6 And thou shalt set the altar of the burnt offering be-

fore the door of the tabernacle of the tent of the congregation.

40:7 And thou shalt set the laver between the tent of the altare holocausti 40:7 labrum inter altare et tabernaculum

congregation and the altar, and shalt put water therein.

40:8 And thou shalt set up the court round about, and hang quod implebis aqua 40:8 circumdabisque atrium tentoriis

up the hanging at the court gate. 40:9 And thou shalt take the anointing oil, and anoint the et ingressum eius 40:9 et adsumpto unctionis oleo ungues tabernacle, and all that is therein, and shalt hallow it, and all the vessels thereof: and it shall be holy.

tabernaculum cum uasis suis ut sanctificentur 40:10 altare

40:10 And thou shalt anoint the altar of the burnt offering, and all his vessels, and sanctify the altar: and it shall be an altar most holy.

holocausti et omnia uasa eius 40:11 labrum cum basi sua

40:11 And thou shalt anoint the laver and his foot, and sanctify it.

omnia unctionis oleo consecrabis ut sint sancta sanctorum

40:12 And thou shalt bring Aaron and his sons unto the 40:12 adplicabisque Aaron et filios eius ad fores tabernaculi

door of the tabernacle of the congregation, and wash them with water.

testimonii et lotos aqua 40:13 indues sanctis uestibus ut mi-

40:13 And thou shalt put upon Aaron the holy garments, and anoint him, and sanctify him; that he may minister unto me in the priest’s office.

nistrent mihi et unctio eorum in sacerdotium proficiat sempi-

40:14 And thou shalt bring his sons, and clothe them with ternum 40:14 fecitque Moses omnia quae praeceperat Docoats:

40:15 And thou shalt anoint them, as thou didst anoint minus 40:15 igitur mense primo anni secundi in prima die

their father, that they may minister unto me in the priest’s office: for their anointing shall surely be an everlasting

priesthood throughout their generations. mensis conlocatum est tabernaculum 40:16 erexitque illud

40:16 Thus did Moses: according to all that the LORD commanded him, so did he.

Moses et posuit tabulas ac bases et uectes statuitque colum-

40:17 And it came to pass in the first month in the second nas 40:17 et expandit tectum super tabernaculum inposito

year, on the first day of the month, that the tabernacle was reared up.

desuper operimento sicut Dominus imperarat 40:18 posuit

40:18 And Moses reared up the tabernacle, and fastened his sockets, and set up the boards thereof, and put in the bars thereof, and reared up his pillars.

et testimonium in arca subditis infra uectibus et oraculum

40:19 And he spread abroad the tent over the taberna- desuper 40:19 cumque intulisset arcam in tabernaculum ad-

cle, and put the covering of the tent above upon it; as the LORD commanded Moses.

pendit ante eam uelum ut expleret Domini iussionem 40:20

40:20 And he took and put the testimony into the ark, and set the staves on the ark, and put the mercy seat above upon the ark:

posuit et mensam in tabernaculo testimonii ad plagam sep-

40:21 And he brought the ark into the tabernacle, and set tentrionalem extra uelum 40:21 ordinatis coram propositio-

up the vail of the covering, and covered the ark of the testimony; as the LORD commanded Moses.

nis panibus sicut praeceperat Dominus Mosi 40:22 posuit

40:22 And he put the table in the tent of the congregation, upon the side of the tabernacle northward, without the vail.

et candelabrum in tabernaculum testimonii e regione men-

40:23 And he set the bread in order upon it before the sae in parte australi 40:23 locatis per ordinem lucernis iuxta

LORD; as the LORD had commanded Moses.

40:24 And he put the candlestick in the tent of the congre- praeceptum Domini 40:24 posuit et altare aureum sub tecto

gation, over against the table, on the side of the tabernacle southward.

testimonii contra uelum 40:25 et adoleuit super eo incen-

40:25 And he lighted the lamps before the LORD; as the LORD commanded Moses.

40:26 And he put the golden altar in the tent of the con- sum aromatum sicut iusserat Dominus 40:26 posuit et ten-

gregation before the vail:

40:27 And he burnt sweet incense thereon; as the LORD torium in introitu tabernaculi 40:27 et altare holocausti in

commanded Moses.

uestibulo testimonii offerens in eo holocaustum et sacrifi-

40:28 And he set up the hanging at the door of the taber- cia ut Dominus imperarat 40:28 labrum quoque statuit inter

nacle.

40:29 And he put the altar of burnt offering by the door of tabernaculum testimonii et altare implens illud aqua 40:29

the tabernacle of the tent of the congregation, and offered upon it the burnt offering and the meat offering; as the

LORD commanded Moses. laueruntque Moses et Aaron ac filii eius manus suas et pe-40:30 And he set the laver between the tent of the congre- des 40:30 cum ingrederentur tectum foederis et accederent

gation and the altar, and put water there, to wash withal.

40:31 And Moses and Aaron and his sons washed their ad altare sicut praeceperat Dominus 40:31 erexit et atrium

hands and their feet thereat:

per gyrum tabernaculi et altaris ducto in introitu eius tentorio postquam cuncta perfecta sunt 40:32 operuit nubes taberna- 40:32 When they went into the tent of the congregation,

and when they came near unto the altar, they washed; as the LORD commanded Moses.

culum testimonii et gloria Domini impleuit illud 40:33 nec

40:33 And he reared up the court round about the tabernacle and the altar, and set up the hanging of the court gate. So Moses finished the work.

poterat Moses ingredi tectum foederis nube operiente omnia et maiestate Domini coruscante quia cuncta nubes operuerat 40:34 si quando nubes tabernaculum deserebat proficis- 40:34 Then a cloud covered the tent of the congregation,

and the glory of the LORD filled the tabernacle.

cebantur filii Israhel per turmas suas 40:35 si pendebat de- 40:35 And Moses was not able to enter into the tent of the

congregation, because the cloud abode thereon, and the glory of the LORD filled the tabernacle.

super manebant in eodem loco 40:36 nubes quippe Domini

40:36 And when the cloud was taken up from over the tabernacle, the children of Israel went onward in all their journeys:

incubabat per diem tabernaculo et ignis in nocte uidentibus populis Israhel per cunctas mansiones suas

Biblia Sacra Vulgata

CLXIV

Liber Leuiticus

1:1 uocauit autem Mosen et locutus est ei Dominus de ta- 1:1 And the LORD called unto Moses, and spake unto him

out of the tabernacle of the congregation, saying,

bernaculo testimonii dicens 1:2 loquere filiis Israhel et di- 1:2 Speak unto the children of Israel, and say unto them, If

any man of you bring an offering unto the LORD, ye shall bring your offering of the cattle, even of the herd, and of

ces ad eos homo qui obtulerit ex uobis hostiam Domino de the flock. pecoribus id est de bubus et ouibus offerens uictimas 1:3 1:3 If his offering be a burnt sacrifice of the herd, let him

offer a male without blemish: he shall offer it of his own voluntary will at the door of the tabernacle of the congre-

si holocaustum fuerit eius oblatio ac de armento masculum gation before the LORD. inmaculatum offeret ad ostium tabernaculi testimonii ad pla-

candum sibi Dominum 1:4 ponetque manus super caput 1:4 And he shall put his hand upon the head of the burnt offering; and it shall be accepted for him to make atonement for him.

hostiae et acceptabilis erit atque in expiationem eius pro-

ficiens 1:5 immolabitque uitulum coram Domino et offe- 1:5 And he shall kill the bullock before the LORD: and the priests, Aaron’s sons, shall bring the blood, and sprinkle the blood round about upon the altar that is by the door of

rent filii Aaron sacerdotes sanguinem eius fundentes super the tabernacle of the congregation. altaris circuitum quod est ante ostium tabernaculi 1:6 de- 1:6 And he shall flay the burnt offering, and cut it into his

pieces.

tractaque pelle hostiae artus in frusta concident 1:7 et su- 1:7 And the sons of Aaron the priest shall put fire upon the

altar, and lay the wood in order upon the fire:

bicient in altari ignem strue lignorum ante conposita 1:8 et 1:8 And the priests, Aaron’s sons, shall lay the parts, the

head, and the fat, in order upon the wood that is on the fire which is upon the altar:

membra quae caesa sunt desuper ordinantes caput uidelicet et cuncta quae adherent iecori 1:9 intestinis et pedibus lo- 1:9 But his inwards and his legs shall he wash in water:

and the priest shall burn all on the altar, to be a burnt sacrifice, an offering made by fire, of a sweet savour unto the

tis aqua adolebitque ea sacerdos super altare in holocaustum LORD. et suauem odorem Domino 1:10 quod si de pecoribus ob- 1:10 And if his offering be of the flocks, namely, of the

sheep, or of the goats, for a burnt sacrifice; he shall bring it a male without blemish.

latio est de ouibus siue de capris holocaustum anniculum et absque macula offeret 1:11 immolabitque ad latus altaris 1:11 And he shall kill it on the side of the altar north-

ward before the LORD: and the priests, Aaron’s sons, shall sprinkle his blood round about upon the altar.

quod respicit ad aquilonem coram Domino sanguinem uero illius fundent super altare filii Aaron per circuitum 1:12 di- 1:12 And he shall cut it into his pieces, with his head and

his fat: and the priest shall lay them in order on the wood that is on the fire which is upon the altar:

uidentque membra caput et omnia quae adherent iecori et inponent super ligna quibus subiciendus est ignis 1:13 inte- 1:13 But he shall wash the inwards and the legs with water:

and the priest shall bring it all, and burn it upon the altar: it is a burnt sacrifice, an offering made by fire, of a sweet

stina uero et pedes lauabunt aqua et oblata omnia adolebit sa- savour unto the LORD. cerdos super altare in holocaustum et odorem suauissimum Domino 1:14 sin autem de auibus holocausti oblatio fuerit 1:14 And if the burnt sacrifice for his offering to the LORD

be of fowls, then he shall bring his offering of turtledoves, or of young pigeons.

Domino de turturibus et pullis columbae 1:15 offeret eam

1:15 And the priest shall bring it unto the altar, and wring off his head, and burn it on the altar; and the blood thereof shall be wrung out at the side of the altar:

sacerdos ad altare et retorto ad collum capite ac rupto uulneris loco decurrere faciet sanguinem super crepidinem altaris 1:16 uesiculam uero gutturis et plumas proiciet propter altare 1:16 And he shall pluck away his crop with his feathers,

and cast it beside the altar on the east part, by the place of the ashes:

ad orientalem plagam in loco in quo cineres effundi solent 1:17 confringetque ascellas eius et non secabit nec ferro di- 1:17 And he shall cleave it with the wings thereof, but

shall not divide it asunder: and the priest shall burn it upon the altar, upon the wood that is upon the fire: it is a burnt sacrifice, an offering made by fire, of a sweet savour unto the LORD.

uidet eam et adolebit super altare lignis igne subposito holo-

Biblia Sacra Vulgata

caustum est et oblatio suauissimi odoris Domino

2:1 And when any will offer a meat offering unto the 2:1 anima cum obtulerit oblationem sacrificii Domino si-

LORD, his offering shall be of fine flour; and he shall pour oil upon it, and put frankincense thereon:

mila erit eius oblatio fundetque super eam oleum et ponet tus

2:2 And he shall bring it to Aaron’s sons the priests: and 2:2 ac deferet ad filios Aaron sacerdotes quorum unus tollet

he shall take thereout his handful of the flour thereof, and of the oil thereof, with all the frankincense thereof; and the priest shall burn the memorial of it upon the altar, to be an offering made by fire, of a sweet savour unto the LORD:

pugillum plenum similae et olei ac totum tus et ponet memo-

2:3 And the remnant of the meat offering shall be Aaron’s riale super altare in odorem suauissimum Domino 2:3 quod

and his sons’: it is a thing most holy of the offerings of the LORD made by fire.

autem reliquum fuerit de sacrificio erit Aaron et filiorum eius

2:4 And if thou bring an oblation of a meat offering baken sanctum sanctorum de oblationibus Domini 2:4 cum autem

in the oven, it shall be unleavened cakes of fine flour mingled with oil, or unleavened wafers anointed with oil.

obtuleris sacrificium coctum in clibano de simila panes scilicet absque fermento conspersos oleo et lagana azyma oleo

2:5 And if thy oblation be a meat offering baken in a pan, lita 2:5 si oblatio tua fuerit de sartagine similae consper-

it shall be of fine flour unleavened, mingled with oil. 2:6 Thou shalt part it in pieces, and pour oil thereon: it is sae oleo et absque fermento 2:6 diuides eam minutatim et a meat offering.

2:7 And if thy oblation be a meat offering baken in the fundes supra oleum 2:7 sin autem de craticula sacrificium

fryingpan, it shall be made of fine flour with oil.

2:8 And thou shalt bring the meat offering that is made of aeque simila oleo conspergetur 2:8 quam offeres Domino

these things unto the LORD: and when it is presented unto the priest, he shall bring it unto the altar.

tradens manibus sacerdotis 2:9 qui cum obtulerit eam tol-

2:9 And the priest shall take from the meat offering a memorial thereof, and shall burn it upon the altar: it is an offering made by fire, of a sweet savour unto the LORD.

let memoriale de sacrificio et adolebit super altare in odo-

2:10 And that which is left of the meat offering shall be rem suauitatis Domino 2:10 quicquid autem reliquum est

Aaron’s and his sons’: it is a thing most holy of the offerings of the LORD made by fire.

erit Aaron et filiorum eius sanctum sanctorum de oblationi-

2:11 No meat offering, which ye shall bring unto the bus Domini 2:11 omnis oblatio quae offertur Domino abs-

LORD, shall be made with leaven: for ye shall burn no leaven, nor any honey, in any offering of the LORD made

by fire. que fermento fiet nec quicquam fermenti ac mellis adole-2:12 As for the oblation of the firstfruits, ye shall offer bitur in sacrificio Domini 2:12 primitias tantum eorum of-

them unto the LORD: but they shall not be burnt on the altar for a sweet savour.

feretis et munera super altare uero non ponentur in odorem

2:13 And every oblation of thy meat offering shalt thou suauitatis 2:13 quicquid obtuleris sacrificii sale condies nec

season with salt; neither shalt thou suffer the salt of the covenant of thy God to be lacking from thy meat offering:

with all thine offerings thou shalt offer salt. auferes sal foederis Dei tui de sacrificio tuo in omni obla-2:14 And if thou offer a meat offering of thy firstfruits tione offeres sal 2:14 sin autem obtuleris munus primarum

unto the LORD, thou shalt offer for the meat offering of thy firstfruits green ears of corn dried by the fire, even corn

beaten out of full ears. frugum tuarum Domino de spicis adhuc uirentibus torres eas

igni et confringes in morem farris et sic offeres primitias tuas

2:15 And thou shalt put oil upon it, and lay frankincense Domino 2:15 fundens supra oleum et tus inponens quia obthereon: it is a meat offering. 2:16 And the priest shall burn the memorial of it, part of latio Domini est 2:16 de qua adolebit sacerdos in memothe beaten corn thereof, and part of the oil thereof, with all the frankincense thereof: it is an offering made by fire

unto the LORD. riam muneris partem farris fracti et olei ac totum tus

3:1 And if his oblation be a sacrifice of peace offering, if 3:1 quod si hostia pacificorum fuerit eius oblatio et de bu-

he offer it of the herd; whether it be a male or female, he shall offer it without blemish before the LORD.

bus uoluerit offerre marem siue feminam inmaculata offe-

3:2 And he shall lay his hand upon the head of his offering, ret coram Domino 3:2 ponetque manum super caput uictiand kill it at the door of the tabernacle of the congregation: and Aaron’s sons the priests shall sprinkle the blood upon the altar round about.

Liber Leuiticus

mae suae quae immolabitur in introitu tabernaculi fundentque filii Aaron sacerdotes sanguinem per circuitum altaris 3:3 et offerent de hostia pacificorum in oblationem Domini 3:3 And he shall offer of the sacrifice of the peace offering

an offering made by fire unto the LORD; the fat that covereth the inwards, and all the fat that is upon the inwards,

adipem qui operit uitalia et quicquid pinguedinis intrinsecus est 3:4 duos renes cum adipe quo teguntur ilia et reticulum 3:4 And the two kidneys, and the fat that is on them, which

is by the flanks, and the caul above the liver, with the kidneys, it shall he take away.

iecoris cum renunculis 3:5 adolebuntque ea super altare in

3:5 And Aaron’s sons shall burn it on the altar upon the burnt sacrifice, which is upon the wood that is on the fire: it is an offering made by fire, of a sweet savour unto the LORD.

holocaustum lignis igne subposito in oblationem suauissimi odoris Domino 3:6 si uero de ouibus fuerit eius oblatio et 3:6 And if his offering for a sacrifice of peace offering

unto the LORD be of the flock; male or female, he shall offer it without blemish.

pacificorum hostia siue masculum siue feminam obtulerit in-

maculata erunt 3:7 si agnum obtulerit coram Domino 3:8 3:7 If he offer a lamb for his offering, then shall he offer it before the LORD. 3:8 And he shall lay his hand upon the head of his offering, and kill it before the tabernacle of the congregation: and Aaron’s sons shall sprinkle the blood thereof round about upon the altar.

ponet manum super caput uictimae suae quae immolabitur in uestibulo tabernaculi testimonii fundentque filii Aaron sanguinem eius per altaris circuitum 3:9 et offerent de pacifi- 3:9 And he shall offer of the sacrifice of the peace offering

an offering made by fire unto the LORD; the fat thereof, and the whole rump, it shall he take off hard by the backbone; and the fat that covereth the inwards, and all the fat that is upon the inwards,

corum hostia sacrificium Domino adipem et caudam totam 3:10 cum renibus et pinguedinem quae operit uentrem atque 3:10 And the two kidneys, and the fat that is upon them,

which is by the flanks, and the caul above the liver, with the kidneys, it shall he take away.

uniuersa uitalia et utrumque renunculum cum adipe qui est iuxta ilia reticulumque iecoris cum renunculis 3:11 et ad- 3:11 And the priest shall burn it upon the altar: it is the

food of the offering made by fire unto the LORD.

olebit ea sacerdos super altare in pabulum ignis et oblationis Domini 3:12 si capra fuerit eius oblatio et obtulerit eam 3:12 And if his offering be a goat, then he shall offer it

before the LORD.

Domino 3:13 ponet manum suam super caput eius immola- 3:13 And he shall lay his hand upon the head of it, and kill it before the tabernacle of the congregation: and the sons of Aaron shall sprinkle the blood thereof upon the altar

bitque eam in introitu tabernaculi testimonii et fundent filii round about. Aaron sanguinem eius per altaris circuitum 3:14 tollentque 3:14 And he shall offer thereof his offering, even an offer-

ing made by fire unto the LORD; the fat that covereth the inwards, and all the fat that is upon the inwards,

ex ea in pastum ignis dominici adipem qui operit uentrem et qui tegit uniuersa uitalia 3:15 duos renunculos cum reticulo 3:15 And the two kidneys, and the fat that is upon them,

which is by the flanks, and the caul above the liver, with the kidneys, it shall he take away.

qui est super eos iuxta ilia et aruinam iecoris cum renunculis 3:16 adolebitque ea sacerdos super altare in alimoniam ig- 3:16 And the priest shall burn them upon the altar: it is the

food of the offering made by fire for a sweet savour: all the fat is the LORD’s.

nis et suauissimi odoris omnis adeps Domini erit 3:17 iure

3:17 It shall be a perpetual statute for your generations throughout all your dwellings, that ye eat neither fat nor blood.

perpetuo in generationibus et cunctis habitaculis uestris nec adipes nec sanguinem omnino comedetis

4:1 locutusque est Dominus ad Mosen dicens 4:2 loquere 4:1 And the LORD spake unto Moses, saying,

4:2 Speak unto the children of Israel, saying, If a soul shall sin through ignorance against any of the commandments of the LORD concerning things which ought not to be done, and shall do against any of them:

filiis Israhel anima cum peccauerit per ignorantiam et de uniuersis mandatis Domini quae praecepit ut non fierent quip-

piam fecerit 4:3 si sacerdos qui est unctus peccauerit de- 4:3 If the priest that is anointed do sin according to the sin of the people; then let him bring for his sin, which he hath sinned, a young bullock without blemish unto the LORD for a sin offering.

Biblia Sacra Vulgata

linquere faciens populum offeret pro peccato suo uitulum

4:4 And he shall bring the bullock unto the door of the inmaculatum Domino 4:4 et adducet illum ad ostium ta-

tabernacle of the congregation before the LORD; and shall lay his hand upon the bullock’s head, and kill the bullock

before the LORD. bernaculi testimonii coram Domino ponetque manum super

4:5 And the priest that is anointed shall take of the bul- caput eius et immolabit eum Domino 4:5 hauriet quoque lock’s blood, and bring it to the tabernacle of the congregation:

de sanguine uituli inferens illud in tabernaculum testimonii

4:6 And the priest shall dip his finger in the blood, and 4:6 cumque intinxerit digitum in sanguinem asperget eo se-

sprinkle of the blood seven times before the LORD, before the vail of the sanctuary.

pties coram Domino contra uelum sanctuarii 4:7 ponetque

4:7 And the priest shall put some of the blood upon the horns of the altar of sweet incense before the LORD, which is in the tabernacle of the congregation; and shall pour all the blood of the bullock at the bottom of the altar of the burnt offering, which is at the door of the tabernacle of the congregation.

de eodem sanguine super cornua altaris thymiamatis gratissimi Domino quod est in tabernaculo testimonii omnem autem reliquum sanguinem fundet in basim altaris holocausti

4:8 And he shall take off from it all the fat of the bullock in introitu tabernaculi 4:8 et adipem uituli auferet pro pec-

for the sin offering; the fat that covereth the inwards, and all the fat that is upon the inwards,

cato tam eum qui operit uitalia quam omnia quae intrinsecus

4:9 And the two kidneys, and the fat that is upon them, sunt 4:9 duos renunculos et reticulum quod est super eos

which is by the flanks, and the caul above the liver, with the kidneys, it shall he take away,

iuxta ilia et adipem iecoris cum renunculis 4:10 sicut aufer-

4:10 As it was taken off from the bullock of the sacrifice of peace offerings: and the priest shall burn them upon the altar of the burnt offering.

tur de uitulo hostiae pacificorum et adolebit ea super altare

4:11 And the skin of the bullock, and all his flesh, with his holocausti 4:11 pellem uero et omnes carnes cum capite et head, and with his legs, and his inwards, and his dung,

4:12 Even the whole bullock shall he carry forth without pedibus et intestinis et fimo 4:12 et reliquo corpore effe-

the camp unto a clean place, where the ashes are poured out, and burn him on the wood with fire: where the ashes

are poured out shall he be burnt. ret extra castra in locum mundum ubi cineres effundi solent

incendetque ea super lignorum struem quae in loco effuso-

4:13 And if the whole congregation of Israel sin through rum cinerum cremabuntur 4:13 quod si omnis turba Israignorance, and the thing be hid from the eyes of the assembly, and they have done somewhat against any of the commandments of the LORD concerning things which should not be done, and are guilty;

hel ignorauerit et per inperitiam fecerit quod contra manda-

4:14 When the sin, which they have sinned against it, is tum Domini est 4:14 et postea intellexerit peccatum suum

known, then the congregation shall offer a young bullock for the sin, and bring him before the tabernacle of the con-

gregation. offeret uitulum pro peccato adducetque eum ad ostium ta-4:15 And the elders of the congregation shall lay their bernaculi 4:15 et ponent seniores populi manus super caput

hands upon the head of the bullock before the LORD: and the bullock shall be killed before the LORD.

eius coram Domino immolatoque uitulo in conspectu Do-

4:16 And the priest that is anointed shall bring of the bul- mini 4:16 inferet sacerdos qui unctus est de sanguine eius in

lock’s blood to the tabernacle of the congregation:

4:17 And the priest shall dip his finger in some of the tabernaculum testimonii 4:17 tincto digito aspergens sep-

blood, and sprinkle it seven times before the LORD, even before the vail.

ties contra uelum 4:18 ponetque de eodem sanguine in cor-

4:18 And he shall put some of the blood upon the horns of the altar which is before the LORD, that is in the tabernacle of the congregation, and shall pour out all the blood at the bottom of the altar of the burnt offering, which is at the door of the tabernacle of the congregation.

nibus altaris quod est coram Domino in tabernaculo testimonii reliquum autem sanguinem fundet iuxta basim altaris

4:19 And he shall take all his fat from him, and burn it holocaustorum quod est in ostio tabernaculi testimonii 4:19

upon the altar.

4:20 And he shall do with the bullock as he did with the omnemque eius adipem tollet et adolebit super altare 4:20

bullock for a sin offering, so shall he do with this: and the priest shall make an atonement for them, and it shall be

forgiven them. sic faciens et de hoc uitulo quomodo fecit et prius et rogante

Liber Leuiticus

pro eis sacerdote propitius erit Dominus 4:21 ipsum autem 4:21 And he shall carry forth the bullock without the

camp, and burn him as he burned the first bullock: it is a sin offering for the congregation.

uitulum efferet extra castra atque conburet sicut et priorem uitulum quia pro peccato est multitudinis 4:22 si peccaue- 4:22 When a ruler hath sinned, and done somewhat

through ignorance against any of the commandments of the LORD his God concerning things which should not be

rit princeps et fecerit unum e pluribus per ignorantiam quod done, and is guilty; Domini lege prohibetur 4:23 et postea intellexerit peccatum 4:23 Or if his sin, wherein he hath sinned, come to his

knowledge; he shall bring his offering, a kid of the goats, a male without blemish:

suum offeret hostiam Domino hircum de capris inmacula-

tum 4:24 ponetque manum suam super caput eius cumque 4:24 And he shall lay his hand upon the head of the goat, and kill it in the place where they kill the burnt offering before the LORD: it is a sin offering.

immolauerit eum in loco ubi solet mactari holocaustum coram Domino quia pro peccato est 4:25 tinguet sacerdos di- 4:25 And the priest shall take of the blood of the sin offer-

ing with his finger, and put it upon the horns of the altar of burnt offering, and shall pour out his blood at the bottom

gitum in sanguine hostiae pro peccato tangens cornua altaris of the altar of burnt offering. holocausti et reliquum fundens ad basim eius 4:26 adipem 4:26 And he shall burn all his fat upon the altar, as the

fat of the sacrifice of peace offerings: and the priest shall make an atonement for him as concerning his sin, and it

uero adolebit supra sicut in uictimis pacificorum fieri solet shall be forgiven him. rogabitque pro eo et pro peccato eius ac dimittetur ei 4:27 4:27 And if any one of the common people sin through ig-

norance, while he doeth somewhat against any of the commandments of the LORD concerning things which ought

quod si peccauerit anima per ignorantiam de populo terrae not to be done, and be guilty; ut faciat quicquam ex his quae Domini lege prohibentur atque delinquat 4:28 et cognouerit peccatum suum offeret ca- 4:28 Or if his sin, which he hath sinned, come to his

knowledge: then he shall bring his offering, a kid of the goats, a female without blemish, for his sin which he hath

pram inmaculatam 4:29 ponetque manum super caput ho- sinned.

4:29 And he shall lay his hand upon the head of the sin offering, and slay the sin offering in the place of the burnt offering.

stiae quae pro peccato est et immolabit eam in loco holocausti 4:30 tolletque sacerdos de sanguine in digito suo et 4:30 And the priest shall take of the blood thereof with

his finger, and put it upon the horns of the altar of burnt offering, and shall pour out all the blood thereof at the

tangens cornua altaris holocausti reliquum fundet ad basim bottom of the altar.

eius 4:31 omnem autem auferens adipem sicut auferri so- 4:31 And he shall take away all the fat thereof, as the fat is taken away from off the sacrifice of peace offerings; and the priest shall burn it upon the altar for a sweet savour unto the LORD; and the priest shall make an atonement for him, and it shall be forgiven him.

let de uictimis pacificorum adolebit super altare in odorem suauitatis Domino rogabitque pro eo et dimittetur ei 4:32 4:32 And if he bring a lamb for a sin offering, he shall

bring it a female without blemish.

sin autem de pecoribus obtulerit uictimam pro peccato ouem scilicet inmaculatam 4:33 ponet manum super caput eius et 4:33 And he shall lay his hand upon the head of the sin

offering, and slay it for a sin offering in the place where they kill the burnt offering.

immolabit eam in loco ubi solent holocaustorum caedi hostiae 4:34 sumetque sacerdos de sanguine eius digito suo et 4:34 And the priest shall take of the blood of the sin offer-

ing with his finger, and put it upon the horns of the altar of burnt offering, and shall pour out all the blood thereof at

tangens cornua altaris holocausti reliquum fundet ad basim the bottom of the altar: eius 4:35 omnem quoque auferens adipem sicut auferri so- 4:35 And he shall take away all the fat thereof, as the fat

of the lamb is taken away from the sacrifice of the peace offerings; and the priest shall burn them upon the altar, according to the offerings made by fire unto the LORD: and the priest shall make an atonement for his sin that he hath committed, and it shall be forgiven him.

let adeps arietis qui immolatur pro pacificis et cremabit super altare in incensum Domini rogabitque pro eo et pro peccato eius et dimittetur illi

5:1 si peccauerit anima et audierit uocem iurantis testis- 5:1 And if a soul sin, and hear the voice of swearing, and

is a witness, whether he hath seen or known of it; if he do not utter it, then he shall bear his iniquity.

que fuerit quod aut ipse uidit aut conscius est nisi indicaue-

5:2 Or if a soul touch any unclean thing, whether it be a rit portabit iniquitatem suam 5:2 anima quae tetigerit ali-

carcase of an unclean beast, or a carcase of unclean cattle, or the carcase of unclean creeping things, and if it be hidden from him; he also shall be unclean, and guilty. quid inmundum siue quod occisum a bestia est aut per se

mortuum uel quodlibet aliud reptile et oblita fuerit inmun-

5:3 Or if he touch the uncleanness of man, whatsoever un- ditiae suae rea est et deliquit 5:3 et si tetigerit quicquam

cleanness it be that a man shall be defiled withal, and it be hid from him; when he knoweth of it, then he shall be

guilty. de inmunditia hominis iuxta omnem inpuritatem qua pol-5:4 Or if a soul swear, pronouncing with his lips to do evil, lui solet oblitaque cognouerit postea subiacebit delicto 5:4

or to do good, whatsoever it be that a man shall pronounce with an oath, and it be hid from him; when he knoweth of

it, then he shall be guilty in one of these. anima quae iurauerit et protulerit labiis suis ut uel male quid faceret uel bene et id ipsum iuramento et sermone firmauerit

5:5 And it shall be, when he shall be guilty in one of these oblitaque postea intellexerit delictum suum 5:5 agat paeni-

things, that he shall confess that he hath sinned in that thing:

tentiam pro peccato 5:6 et offerat agnam de gregibus siue

5:6 And he shall bring his trespass offering unto the LORD for his sin which he hath sinned, a female from the flock, a lamb or a kid of the goats, for a sin offering; and the priest shall make an atonement for him concerning his sin.

capram orabitque pro eo sacerdos et pro peccato eius 5:7

5:7 And if he be not able to bring a lamb, then he shall bring for his trespass, which he hath committed, two turtledoves, or two young pigeons, unto the LORD; one for a sin offering, and the other for a burnt offering.

sin autem non potuerit offerre pecus offerat duos turtures uel duos pullos columbarum Domino unum pro peccato et alte-

5:8 And he shall bring them unto the priest, who shall offer rum in holocaustum 5:8 dabitque eos sacerdoti qui primum that which is for the sin offering first, and wring off his head from his neck, but shall not divide it asunder:

offerens pro peccato retorquebit caput eius ad pinnulas ita ut

5:9 And he shall sprinkle of the blood of the sin offering collo hereat et non penitus abrumpatur 5:9 et asperget de

upon the side of the altar; and the rest of the blood shall be wrung out at the bottom of the altar: it is a sin offering.

sanguine eius parietem altaris quicquid autem reliquum fuerit faciet destillare ad fundamentum eius quia pro peccato

5:10 And he shall offer the second for a burnt offering, ac- est 5:10 alterum uero adolebit holocaustum ut fieri solet ro-

cording to the manner: and the priest shall make an atonement for him for his sin which he hath sinned, and it shall

be forgiven him. gabitque pro eo sacerdos et pro peccato eius et dimittetur

5:11 But if he be not able to bring two turtledoves, or two ei 5:11 quod si non quiuerit manus eius offerre duos turyoung pigeons, then he that sinned shall bring for his offering the tenth part of an ephah of fine flour for a sin offering; he shall put no oil upon it, neither shall he put any frankincense thereon: for it is a sin offering.

tures uel duos pullos columbae offeret pro peccato similam partem oephi decimam non mittet in eam oleum nec turis ali-

5:12 Then shall he bring it to the priest, and the priest shall quid inponet quia pro peccato est 5:12 tradetque eam sacertake his handful of it, even a memorial thereof, and burn it on the altar, according to the offerings made by fire unto

the LORD: it is a sin offering. doti qui plenum ex toto pugillum hauriens cremabit super al-5:13 And the priest shall make an atonement for him as tare in monumentum eius qui obtulit 5:13 rogans pro illo et

touching his sin that he hath sinned in one of these, and it shall be forgiven him: and the remnant shall be the priest’s,

as a meat offering. expians reliquam uero partem ipse habebit in munere 5:14

5:14 And the LORD spake unto Moses, saying,

5:15 If a soul commit a trespass, and sin through igno- locutus est Dominus ad Mosen dicens 5:15 anima si prae-

rance, in the holy things of the LORD; then he shall bring for his trespass unto the LORD a ram without blemish out of the flocks, with thy estimation by shekels of silver, after the shekel of the sanctuary, for a trespass offering.

uaricans caerimonias per errorem in his quae Domino sunt sanctificata peccauerit offeret pro delicto suo arietem inmaculatum de gregibus qui emi potest duobus siclis iuxta pon-

5:16 And he shall make amends for the harm that he hath dus sanctuarii 5:16 ipsumque quod intulit damni restituet

done in the holy thing, and shall add the fifth part thereto, and give it unto the priest: and the priest shall make an atonement for him with the ram of the trespass offering, and it shall be forgiven him.

Liber Leuiticus

et quintam partem ponet supra tradens sacerdoti qui rogabit pro eo offerens arietem et dimittetur ei 5:17 anima si pec- 5:17 And if a soul sin, and commit any of these things

which are forbidden to be done by the commandments of the LORD; though he wist it not, yet is he guilty, and shall

cauerit per ignorantiam feceritque unum ex his quae Domini bear his iniquity. lege prohibentur et peccati rea intellexerit iniquitatem suam 5:18 offeret arietem inmaculatum de gregibus sacerdoti iuxta 5:18 And he shall bring a ram without blemish out of the

flock, with thy estimation, for a trespass offering, unto the priest: and the priest shall make an atonement for him concerning his ignorance wherein he erred and wist it not, and it shall be forgiven him.

mensuram aestimationemque peccati qui orabit pro eo quod nesciens fecerit et dimittetur ei 5:19 quia per errorem deli- 5:19 It is a trespass offering: he hath certainly trespassed

against the LORD.

quit in Dominum

6:1 locutus est Dominus ad Mosen dicens 6:2 anima quae 6:1 And the LORD spake unto Moses, saying,

6:2 If a soul sin, and commit a trespass against the LORD, and lie unto his neighbour in that which was delivered him to keep, or in fellowship, or in a thing taken away by violence, or hath deceived his neighbour;

peccauerit et contempto Domino negauerit depositum proximo suo quod fidei eius creditum fuerat uel ui aliquid extorserit aut calumniam fecerit 6:3 siue rem perditam inue- 6:3 Or have found that which was lost, and lieth concern-

ing it, and sweareth falsely; in any of all these that a man doeth, sinning therein:

nerit et infitians insuper peierarit et quodlibet aliud ex pluribus fecerit in quibus peccare solent homines 6:4 conuicta 6:4 Then it shall be, because he hath sinned, and is guilty,

that he shall restore that which he took violently away, or the thing which he hath deceitfully gotten, or that which was delivered him to keep, or the lost thing which he found,

delicti reddet 6:5 omnia quae per fraudem uoluit obtinere

6:5 Or all that about which he hath sworn falsely; he shall even restore it in the principal, and shall add the fifth part more thereto, and give it unto him to whom it appertaineth, in the day of his trespass offering.

integra et quintam insuper partem domino cui damnum intulerat 6:6 pro peccato autem suo offeret arietem inmacu-

6:6 And he shall bring his trespass offering unto the LORD, a ram without blemish out of the flock, with thy

latum de grege et dabit eum sacerdoti iuxta aestimationem _{estimation, for a trespass offering, unto the priest:} mensuramque delicti 6:7 qui rogabit pro eo coram Domino 6:7 And the priest shall make an atonement for him before

the LORD: and it shall be forgiven him for any thing of all that he hath done in trespassing therein.

et dimittetur illi pro singulis quae faciendo peccauerit 6:8

6:8 And the LORD spake unto Moses, saying,

locutus est Dominus ad Mosen dicens 6:9 praecipe Aaron 6:9 Command Aaron and his sons, saying, This is the law

of the burnt offering: It is the burnt offering, because of the burning upon the altar all night unto the morning, and

et filiis eius haec est lex holocausti cremabitur in altari tota the fire of the altar shall be burning in it. nocte usque mane ignis ex eodem altari erit 6:10 uestietur 6:10 And the priest shall put on his linen garment, and his

linen breeches shall he put upon his flesh, and take up the ashes which the fire hath consumed with the burnt offering

sacerdos tunica et feminalibus lineis tolletque cineres quos on the altar, and he shall put them beside the altar. uorans ignis exusit et ponens iuxta altare 6:11 spoliabitur 6:11 And he shall put off his garments, and put on other

garments, and carry forth the ashes without the camp unto a clean place.

prioribus uestimentis indutusque aliis efferet eos extra castra et in loco mundissimo usque ad fauillam consumi faciet 6:12 ignis autem in altari semper ardebit quem nutriet sacer- 6:12 And the fire upon the altar shall be burning in it; it

shall not be put out: and the priest shall burn wood on it every morning, and lay the burnt offering in order upon it;

dos subiciens ligna mane per singulos dies et inposito ho- and he shall burn thereon the fat of the peace offerings. locausto desuper adolebit adipes pacificorum 6:13 ignis est 6:13 The fire shall ever be burning upon the altar; it shall

never go out.

iste perpetuus qui numquam deficiet in altari 6:14 haec est 6:14 And this is the law of the meat offering: the sons of

Aaron shall offer it before the LORD, before the altar.

lex sacrificii et libamentorum quae offerent filii Aaron coram Domino et coram altari 6:15 tollet sacerdos pugillum simi- 6:15 And he shall take of it his handful, of the flour of the

meat offering, and of the oil thereof, and all the frankincense which is upon the meat offering, and shall burn it upon the altar for a sweet savour, even the memorial of it, unto the LORD.

Biblia Sacra Vulgata

lae quae conspersa est oleo et totum tus quod super similam positum est adolebitque illud in altari in monumentum odo-

6:16 And the remainder thereof shall Aaron and his sons ris suauissimi Domino 6:16 reliquam autem partem similae

eat: with unleavened bread shall it be eaten in the holy place; in the court of the tabernacle of the congregation

they shall eat it. comedet Aaron cum filiis suis absque fermento et comedet 6:17 It shall not be baken with leaven. I have given it in loco sancto atrii tabernaculi 6:17 ideo autem non fermen-

unto them for their portion of my offerings made by fire; it is most holy, as is the sin offering, and as the trespass

offering. tabitur quia pars eius in Domini offertur incensum sanctum 6:18 All the males among the children of Aaron shall eat sanctorum erit sicut pro peccato atque delicto 6:18 mares

of it. It shall be a statute for ever in your generations concerning the offerings of the LORD made by fire: every one

that toucheth them shall be holy. tantum stirpis Aaron comedent illud legitimum ac sempiter-

num est in generationibus uestris de sacrificiis Domini omnis

6:19 And the LORD spake unto Moses, saying, qui tetigerit illa sanctificabitur 6:19 et locutus est Dominus

6:20 This is the offering of Aaron and of his sons, which ad Mosen dicens 6:20 haec est oblatio Aaron et filiorum they shall offer unto the LORD in the day when he is anointed; the tenth part of an ephah of fine flour for a meat offering perpetual, half of it in the morning, and half thereof at night.

eius quam offerre debent Domino in die unctionis suae decimam partem oephi offerent similae in sacrificio sempiterno

6:21 In a pan it shall be made with oil; and when it is medium eius mane et medium uespere 6:21 quae in sarta-

baken, thou shalt bring it in: and the baken pieces of the meat offering shalt thou offer for a sweet savour unto the

LORD. gine oleo conspersa frigetur offeret autem eam calidam in

6:22 And the priest of his sons that is anointed in his stead odorem suauissimum Domino 6:22 sacerdos qui patri iure shall offer it: it is a statute for ever unto the LORD; it shall be wholly burnt.

successerit et tota cremabitur in altari 6:23 omne enim sa-

6:23 For every meat offering for the priest shall be wholly burnt: it shall not be eaten.

crificium sacerdotum igne consumetur nec quisquam come-

6:24 And the LORD spake unto Moses, saying, det ex eo 6:24 locutus est Dominus ad Mosen dicens 6:25

6:25 Speak unto Aaron and to his sons, saying, This is the law of the sin offering: In the place where the burnt offering is killed shall the sin offering be killed before the LORD: it is most holy.

loquere Aaron et filiis eius ista est lex hostiae pro peccato in loco ubi offertur holocaustum immolabitur coram Domino

6:26 The priest that offereth it for sin shall eat it: in the sanctum sanctorum est 6:26 sacerdos qui offert comedet holy place shall it be eaten, in the court of the tabernacle of the congregation.

eam in loco sancto in atrio tabernaculi 6:27 quicquid teti-

6:27 Whatsoever shall touch the flesh thereof shall be holy: and when there is sprinkled of the blood thereof upon any garment, thou shalt wash that whereon it was sprinkled in the holy place.

gerit carnes eius sanctificabitur si de sanguine illius uestis

6:28 But the earthen vessel wherein it is sodden shall be fuerit aspersa lauabitur in loco sancto 6:28 uas autem fic-

broken: and if it be sodden in a brasen pot, it shall be both scoured, and rinsed in water.

tile in quo cocta est confringetur quod si uas aeneum fuerit

6:29 All the males among the priests shall eat thereof: it is defricabitur et lauabitur aqua 6:29 omnis masculus de gemost holy.

nere sacerdotali uescetur carnibus eius quia sanctum sancto-

6:30 And no sin offering, whereof any of the blood is rum est 6:30 hostia enim quae caeditur pro peccato cuius

brought into the tabernacle of the congregation to reconcile withal in the holy place, shall be eaten: it shall be

burnt in the fire. sanguis infertur in tabernaculum testimonii ad expiandum in sanctuario non comedetur sed conburetur igni

7:1 Likewise this is the law of the trespass offering: it is 7:1 haec quoque est lex hostiae pro delicto sancta sancmost holy.

7:2 In the place where they kill the burnt offering shall torum est 7:2 idcirco ubi immolatur holocaustum macta-

they kill the trespass offering: and the blood thereof shall he sprinkle round about upon the altar.

Liber Leuiticus

bitur et uictima pro delicto sanguis eius per gyrum fundetur altaris 7:3 offerent ex ea caudam et adipem qui ope- 7:3 And he shall offer of it all the fat thereof; the rump,

and the fat that covereth the inwards,

rit uitalia 7:4 duos renunculos et pinguedinem quae iuxta 7:4 And the two kidneys, and the fat that is on them, which

is by the flanks, and the caul that is above the liver, with the kidneys, it shall he take away:

ilia est reticulumque iecoris cum renunculis 7:5 et adole-

7:5 And the priest shall burn them upon the altar for an offering made by fire unto the LORD: it is a trespass offering.

bit ea sacerdos super altare incensum est Domini pro delicto 7:6 omnis masculus de sacerdotali genere in loco sancto ue- 7:6 Every male among the priests shall eat thereof: it shall

be eaten in the holy place: it is most holy.

scetur his carnibus quia sanctum sanctorum est 7:7 sicut 7:7 As the sin offering is, so is the trespass offering: there is one law for them: the priest that maketh atonement therewith shall have it.

pro peccato offertur hostia ita et pro delicto utriusque hostiae lex una erit ad sacerdotem qui eam obtulerit pertinebit 7:8 sacerdos qui offert holocausti uictimam habebit pellem 7:8 And the priest that offereth any man’s burnt offering,

even the priest shall have to himself the skin of the burnt offering which he hath offered.

eius 7:9 et omne sacrificium similae quod coquitur in cli-

7:9 And all the meat offering that is baken in the oven, and all that is dressed in the fryingpan, and in the pan, shall be the priest’s that offereth it.

bano et quicquid in craticula uel in sartagine praeparatur eius erit sacerdotis a quo offertur 7:10 siue oleo conspersa siue 7:10 And every meat offering, mingled with oil, and dry,

shall all the sons of Aaron have, one as much as another.

arida fuerit cunctis filiis Aaron aequa mensura per singulos diuidetur 7:11 haec est lex hostiae pacificorum quae of- 7:11 And this is the law of the sacrifice of peace offerings,

which he shall offer unto the LORD.

fertur Domino 7:12 si pro gratiarum actione fuerit oblatio 7:12 If he offer it for a thanksgiving, then he shall offer

with the sacrifice of thanksgiving unleavened cakes mingled with oil, and unleavened wafers anointed with oil,

offerent panes absque fermento conspersos oleo et lagana and cakes mingled with oil, of fine flour, fried. azyma uncta oleo coctamque similam et collyridas olei ad-

mixtione conspersas 7:13 panes quoque fermentatos cum 7:13 Besides the cakes, he shall offer for his offering leavened bread with the sacrifice of thanksgiving of his peace offerings.

hostia gratiarum quae immolatur pro pacificis 7:14 ex qui-

7:14 And of it he shall offer one out of the whole oblation for an heave offering unto the LORD, and it shall be the priest’s that sprinkleth the blood of the peace offerings.

bus unus pro primitiis offeretur Domino et erit sacerdotis qui

fundet hostiae sanguinem 7:15 cuius carnes eadem come- 7:15 And the flesh of the sacrifice of his peace offerings for thanksgiving shall be eaten the same day that it is offered; he shall not leave any of it until the morning.

dentur die nec remanebit ex eis quicquam usque mane 7:16

7:16 But if the sacrifice of his offering be a vow, or a voluntary offering, it shall be eaten the same day that he offereth his sacrifice: and on the morrow also the remainder of it shall be eaten:

si uoto uel sponte quisquam obtulerit hostiam eadem similiter edetur die sed et si quid in crastinum remanserit uesci licitum est 7:17 quicquid autem tertius inuenerit dies ignis 7:17 But the remainder of the flesh of the sacrifice on the

third day shall be burnt with fire.

absumet 7:18 si quis de carnibus uictimae pacificorum die 7:18 And if any of the flesh of the sacrifice of his peace

offerings be eaten at all on the third day, it shall not be accepted, neither shall it be imputed unto him that offereth it: it shall be an abomination, and the soul that eateth of it shall bear his iniquity.

tertio comederit irrita fiet oblatio nec prodeeit offerenti quin potius quaecumque anima tali se edulio contaminarit praeuaricationis rea erit 7:19 caro quae aliquid tetigerit inmundum 7:19 And the flesh that toucheth any unclean thing shall

not be eaten; it shall be burnt with fire: and as for the flesh, all that be clean shall eat thereof.

non comedetur sed conburetur igni qui fuerit mundus uescetur ea 7:20 anima polluta quae ederit de carnibus hostiae 7:20 But the soul that eateth of the flesh of the sacrifice

of peace offerings, that pertain unto the LORD, having his uncleanness upon him, even that soul shall be cut off from

pacificorum quae oblata est Domino peribit de populis suis his people.

Biblia Sacra Vulgata

7:21 Moreover the soul that shall touch any unclean thing, 7:21 et quae tetigerit inmunditiam hominis uel iumenti siue

as the uncleanness of man, or any unclean beast, or any abominable unclean thing, and eat of the flesh of the sacrifice of peace offerings, which pertain unto the LORD, even that soul shall be cut off from his people.

omnis rei quae polluere potest et comederit de huiuscemodi

7:22 And the LORD spake unto Moses, saying, carnibus interibit de populis suis 7:22 locutusque est Do-

7:23 Speak unto the children of Israel, saying, Ye shall eat minus ad Mosen dicens 7:23 loquere filiis Israhel adipem no manner of fat, of ox, or of sheep, or of goat. 7:24 And the fat of the beast that dieth of itself, and the fat bouis et ouis et caprae non comedetis 7:24 adipem cadaueof that which is torn with beasts, may be used in any other use: but ye shall in no wise eat of it.

ris morticini et eius animalis quod a bestia captum est ha-

7:25 For whosoever eateth the fat of the beast, of which bebitis in usus uarios 7:25 si quis adipem qui offerri debet

men offer an offering made by fire unto the LORD, even the soul that eateth it shall be cut off from his people.

in incensum Domini comederit peribit de populo suo 7:26

7:26 Moreover ye shall eat no manner of blood, whether it be of fowl or of beast, in any of your dwellings.

sanguinem quoque omnis animalis non sumetis in cibo tam

7:27 Whatsoever soul it be that eateth any manner of de auibus quam de pecoribus 7:27 omnis anima quae ederit

blood, even that soul shall be cut off from his people.

7:28 And the LORD spake unto Moses, saying, sanguinem peribit de populis suis 7:28 locutus est Dominus 7:29 Speak unto the children of Israel, saying, He that of- ad Mosen dicens 7:29 loquere filiis Israhel qui offert uicti-

fereth the sacrifice of his peace offerings unto the LORD shall bring his oblation unto the LORD of the sacrifice of

his peace offerings. mam pacificorum Domino offerat simul et sacrificium id est 7:30 His own hands shall bring the offerings of the LORD libamenta eius 7:30 tenebit manibus adipem hostiae et pec-

made by fire, the fat with the breast, it shall he bring, that the breast may be waved for a wave offering before the

LORD. tusculum cumque ambo oblata Domino consecrarit tradet sa-7:31 And the priest shall burn the fat upon the altar: but cerdoti 7:31 qui adolebit adipem super altare pectusculum

the breast shall be Aaron’s and his sons’. 7:32 And the right shoulder shall ye give unto the priest for autem erit Aaron et filiorum eius 7:32 armus quoque dexter an heave offering of the sacrifices of your peace offerings.

7:33 He among the sons of Aaron, that offereth the blood de pacificorum hostiis cedet in primitias sacerdotis 7:33 qui

of the peace offerings, and the fat, shall have the right shoulder for his part.

obtulerit sanguinem et adipem filiorum Aaron ipse habebit

7:34 For the wave breast and the heave shoulder have I et armum dextrum in portione sua 7:34 pectusculum enim

taken of the children of Israel from off the sacrifices of their peace offerings, and have given them unto Aaron the priest and unto his sons by a statute for ever from among the children of Israel.

elationis et armum separationis tuli a filiis Israhel de hostiis eorum pacificis et dedi Aaron sacerdoti ac filiis eius lege per-

7:35 This is the portion of the anointing of Aaron, and petua ab omni populo Israhel 7:35 haec est unctio Aaron et

of the anointing of his sons, out of the offerings of the LORD made by fire, in the day when he presented them to

minister unto the LORD in the priest’s office; filiorum eius in caerimoniis Domini die qua obtulit eos Mo-7:36 Which the LORD commanded to be given them of ses ut sacerdotio fungerentur 7:36 et quae praecepit dari eis

the children of Israel, in the day that he anointed them, by a statute for ever throughout their generations.

Dominus a filiis Israhel religione perpetua in generationi-

7:37 This is the law of the burnt offering, of the meat of- bus suis 7:37 ista est lex holocausti et sacrificii pro peccato

fering, and of the sin offering, and of the trespass offering, and of the consecrations, and of the sacrifice of the peace

offerings; atque delicto et pro consecratione et pacificorum uictimis 7:38 Which the LORD commanded Moses in mount Sinai, 7:38 quas constituit Dominus Mosi in monte Sinai quando

in the day that he commanded the children of Israel to offer their oblations unto the LORD, in the wilderness of Sinai.

mandauit filiis Israhel ut offerrent oblationes suas Domino in deserto Sinai

8:1 And the LORD spake unto Moses, saying, 8:1 locutusque est Dominus ad Mosen dicens 8:2 tolle

8:2 Take Aaron and his sons with him, and the garments, and the anointing oil, and a bullock for the sin offering, and two rams, and a basket of unleavened bread;

Aaron cum filiis suis uestes eorum et unctionis oleum uitu-

Liber Leuiticus

lum pro peccato duos arietes canistrum cum azymis 8:3 et 8:3 And gather thou all the congregation together unto the

door of the tabernacle of the congregation.

congregabis omnem coetum ad ostium tabernaculi 8:4 fe- 8:4 And Moses did as the LORD commanded him; and

the assembly was gathered together unto the door of the tabernacle of the congregation.

cit Moses ut Dominus imperarat congregataque omni turba ante fores 8:5 ait iste est sermo quem iussit Dominus fieri 8:5 And Moses said unto the congregation, This is the

thing which the LORD commanded to be done.

8:6 statimque obtulit Aaron et filios eius cumque lauisset eos 8:6 And Moses brought Aaron and his sons, and washed

them with water.

8:7 uestiuit pontificem subucula linea accingens eum balteo 8:7 And he put upon him the coat, and girded him with the

girdle, and clothed him with the robe, and put the ephod upon him, and he girded him with the curious girdle of the

et induens tunica hyacinthina et desuper umerale inposuit ephod, and bound it unto him therewith. 8:8 quod adstringens cingulo aptauit rationali in quo erat doc- 8:8 And he put the breastplate upon him: also he put in the

breastplate the Urim and the Thummim.

trina et ueritas 8:9 cidarim quoque texit caput et super eam 8:9 And he put the mitre upon his head; also upon the

mitre, even upon his forefront, did he put the golden plate, the holy crown; as the LORD commanded Moses.

contra frontem posuit lamminam auream consecratam in sanctificationem sicut praeceperat ei Dominus 8:10 tulit et 8:10 And Moses took the anointing oil, and anointed the

tabernacle and all that was therein, and sanctified them.

unctionis oleum quo leuit tabernaculum cum omni supellectili sua 8:11 cumque sanctificans aspersisset altare septem 8:11 And he sprinkled thereof upon the altar seven times,

and anointed the altar and all his vessels, both the laver and his foot, to sanctify them.

uicibus unxit illud et omnia uasa eius labrumque cum basi sua sanctificauit oleo 8:12 quod fundens super caput Aaron 8:12 And he poured of the anointing oil upon Aaron’s

head, and anointed him, to sanctify him.

unxit eum et consecrauit 8:13 filios quoque eius oblatos ue- 8:13 And Moses brought Aaron’s sons, and put coats upon

them, and girded them with girdles, and put bonnets upon them; as the LORD commanded Moses.

stiuit tunicis lineis et cinxit balteo inposuitque mitras ut iusserat Dominus 8:14 obtulit et uitulum pro peccato cumque 8:14 And he brought the bullock for the sin offering: and

Aaron and his sons laid their hands upon the head of the bullock for the sin offering.

super caput eius posuissent Aaron et filii eius manus suas 8:15 immolauit eum hauriens sanguinem et tincto digito teti- 8:15 And he slew it; and Moses took the blood, and put

it upon the horns of the altar round about with his finger, and purified the altar, and poured the blood at the bottom of the altar, and sanctified it, to make reconciliation upon it.

git cornua altaris per gyrum quo expiato et sanctificato fudit

reliquum sanguinem ad fundamenta eius 8:16 adipem au- 8:16 And he took all the fat that was upon the inwards, and the caul above the liver, and the two kidneys, and their fat, and Moses burned it upon the altar.

tem qui erat super uitalia et reticulum iecoris duosque renunculos cum aruinulis suis adoleuit super altare 8:17 uitulum 8:17 But the bullock, and his hide, his flesh, and his dung,

he burnt with fire without the camp; as the LORD commanded Moses.

cum pelle carnibus et fimo cremans extra castra sicut praeceperat Dominus 8:18 obtulit et arietem in holocaustum su- 8:18 And he brought the ram for the burnt offering: and

Aaron and his sons laid their hands upon the head of the ram.

per cuius caput cum inposuissent Aaron et filii eius manus suas 8:19 immolauit eum et fudit sanguinem eius per altaris 8:19 And he killed it; and Moses sprinkled the blood upon

the altar round about.

circuitum 8:20 ipsumque arietem in frusta concidens caput 8:20 And he cut the ram into pieces; and Moses burnt the

head, and the pieces, and the fat.

eius et artus et adipem adoleuit igni 8:21 lotis prius intesti- 8:21 And he washed the inwards and the legs in water; and

Moses burnt the whole ram upon the altar: it was a burnt sacrifice for a sweet savour, and an offering made by fire

nis et pedibus totumque simul arietem incendit super altare unto the LORD; as the LORD commanded Moses. eo quod esset holocaustum suauissimi odoris Domino sicut praeceperat ei 8:22 obtulit et arietem secundum in consec- 8:22 And he brought the other ram, the ram of consecra-

tion: and Aaron and his sons laid their hands upon the head of the ram.

Biblia Sacra Vulgata

rationem sacerdotum posueruntque super caput illius Aaron

8:23 And he slew it; and Moses took of the blood of it, et filii eius manus suas 8:23 quem cum immolasset Moses

and put it upon the tip of Aaron’s right ear, and upon the thumb of his right hand, and upon the great toe of his right

foot. sumens de sanguine tetigit extremum auriculae dextrae Aa-8:24 And he brought Aaron’s sons, and Moses put of the ron et pollicem manus eius dextrae similiter et pedis 8:24

blood upon the tip of their right ear, and upon the thumbs of their right hands, and upon the great toes of their right feet: and Moses sprinkled the blood upon the altar round about.

obtulit et filios Aaron cumque de sanguine arietis immolati tetigisset extremum auriculae singulorum dextrae et pollices manus ac pedis dextri reliquum fudit super altare per cir-

8:25 And he took the fat, and the rump, and all the fat that cuitum 8:25 adipem uero et caudam omnemque pinguediwas upon the inwards, and the caul above the liver, and the two kidneys, and their fat, and the right shoulder:

nem quae operit intestina reticulumque iecoris et duos renes

8:26 And out of the basket of unleavened bread, that was cum adipibus suis et armo dextro separauit 8:26 tollens au-

before the LORD, he took one unleavened cake, and a cake of oiled bread, and one wafer, and put them on the

fat, and upon the right shoulder: tem de canistro azymorum quod erat coram Domino panem

absque fermento et collyridam conspersam oleo laganumque

8:27 And he put all upon Aaron’s hands, and upon his posuit super adipes et armum dextrum 8:27 tradens simul

sons’ hands, and waved them for a wave offering before the LORD.

omnia Aaron et filiis eius qui postquam leuauerunt ea co-

8:28 And Moses took them from off their hands, and burnt ram Domino 8:28 rursum suscepta de manibus eorum adthem on the altar upon the burnt offering: they were consecrations for a sweet savour: it is an offering made by fire

unto the LORD. oleuit super altare holocausti eo quod consecrationis esset 8:29 And Moses took the breast, and waved it for a wave oblatio in odorem suauitatis sacrificii Domini 8:29 tulit et

offering before the LORD: for of the ram of consecration it was Moses’ part; as the LORD commanded Moses.

pectusculum eleuans illud coram Domino de ariete consec-

8:30 And Moses took of the anointing oil, and of the blood rationis in partem suam sicut praeceperat ei Dominus 8:30

which was upon the altar, and sprinkled it upon Aaron, and upon his garments, and upon his sons, and upon his sons’ garments with him; and sanctified Aaron, and his garments, and his sons, and his sons’ garments with him.

adsumensque unguentum et sanguinem qui erat in altari aspersit super Aaron et uestimenta eius et super filios illius ac

8:31 And Moses said unto Aaron and to his sons, Boil the uestes eorum 8:31 cumque sanctificasset eos in uestitu suo

flesh at the door of the tabernacle of the congregation: and there eat it with the bread that is in the basket of consecrations, as I commanded, saying, Aaron and his sons shall eat it.

praecepit eis dicens coquite carnes ante fores tabernaculi et ibi comedite eas panes quoque consecrationis edite qui positi sunt in canistro sicut praecepit mihi dicens Aaron et filii

8:32 And that which remaineth of the flesh and of the eius comedent eos 8:32 quicquid autem reliquum fuerit de

bread shall ye burn with fire.

8:33 And ye shall not go out of the door of the tabernacle carne et panibus ignis absumet 8:33 de ostio quoque taber-

of the congregation in seven days, until the days of your consecration be at an end: for seven days shall he conse-

crate you. naculi non exibitis septem diebus usque ad diem quo con-

plebitur tempus consecrationis uestrae septem enim diebus

8:34 As he hath done this day, so the LORD hath com- finitur consecratio 8:34 sicut et inpraesentiarum factum est

manded to do, to make an atonement for you.

8:35 Therefore shall ye abide at the door of the tabernacle ut ritus sacrificii conpleretur 8:35 die ac nocte manebitis in

of the congregation day and night seven days, and keep the charge of the LORD, that ye die not: for so I am com-

manded. tabernaculo obseruantes custodias Domini ne moriamini sic 8:36 So Aaron and his sons did all things which the LORD enim mihi praeceptum est 8:36 feceruntque Aaron et filii

commanded by the hand of Moses.

Liber Leuiticus

eius cuncta quae locutus est Dominus per manum Mosi

9:1 facto autem octauo die uocauit Moses Aaron et filios 9:1 And it came to pass on the eighth day, that Moses

called Aaron and his sons, and the elders of Israel;

eius ac maiores natu Israhel dixitque ad Aaron 9:2 tolle de 9:2 And he said unto Aaron, Take thee a young calf for a

sin offering, and a ram for a burnt offering, without blemish, and offer them before the LORD.

armento uitulum pro peccato et arietem in holocaustum ut-

rumque inmaculatos et offer illos coram Domino 9:3 et ad 9:3 And unto the children of Israel thou shalt speak, saying, Take ye a kid of the goats for a sin offering; and a calf and a lamb, both of the first year, without blemish, for a

filios Israhel loqueris tollite hircum pro peccato et uitulum burnt offering; atque agnum anniculos et sine macula in holocaustum 9:4 9:4 Also a bullock and a ram for peace offerings, to sac-

rifice before the LORD; and a meat offering mingled with oil: for to day the LORD will appear unto you.

bouem et arietem pro pacificis et immolate eos coram Domino in sacrificio singulorum similam oleo conspersam offerentes hodie enim Dominus apparebit uobis 9:5 tulerunt 9:5 And they brought that which Moses commanded be-

fore the tabernacle of the congregation: and all the congregation drew near and stood before the LORD.

ergo cuncta quae iusserat Moses ad ostium tabernaculi ubi cum omnis staret multitudo 9:6 ait Moses iste est sermo 9:6 And Moses said, This is the thing which the LORD

commanded that ye should do: and the glory of the LORD shall appear unto you.

quem praecepit Dominus facite et apparebit uobis gloria eius 9:7 dixit et ad Aaron accede ad altare et immola pro peccato 9:7 And Moses said unto Aaron, Go unto the altar, and

offer thy sin offering, and thy burnt offering, and make an atonement for thyself, and for the people: and offer the offering of the people, and make an atonement for them; as the LORD commanded.

tuo offer holocaustum et deprecare pro te et pro populo cumque mactaueris hostiam populi ora pro eo sicut praecepit Do-

minus 9:8 statimque Aaron accedens ad altare immolauit 9:8 Aaron therefore went unto the altar, and slew the calf

of the sin offering, which was for himself.

uitulum pro peccato suo 9:9 cuius sanguinem obtulerunt ei 9:9 And the sons of Aaron brought the blood unto him:

and he dipped his finger in the blood, and put it upon the horns of the altar, and poured out the blood at the bottom

filii sui in quo tinguens digitum tetigit cornua altaris et fudit of the altar: residuum ad basim eius 9:10 adipemque et renunculos ac 9:10 But the fat, and the kidneys, and the caul above the

liver of the sin offering, he burnt upon the altar; as the LORD commanded Moses.

reticulum iecoris quae sunt pro peccato adoleuit super altare sicut praeceperat Dominus Mosi 9:11 carnes uero et pellem 9:11 And the flesh and the hide he burnt with fire without

the camp.

eius extra castra conbusit igni 9:12 immolauit et holocausti 9:12 And he slew the burnt offering; and Aaron’s sons pre-

sented unto him the blood, which he sprinkled round about upon the altar.

uictimam obtuleruntque ei filii sui sanguinem eius quem fudit per altaris circuitum 9:13 ipsam etiam hostiam in frusta 9:13 And they presented the burnt offering unto him, with

the pieces thereof, and the head: and he burnt them upon the altar.

concisam cum capite et membris singulis obtulerunt quae omnia super altare cremauit igni 9:14 lotis prius aqua in- 9:14 And he did wash the inwards and the legs, and burnt

them upon the burnt offering on the altar.

testinis et pedibus 9:15 et pro peccato populi offerens mac- 9:15 And he brought the people’s offering, and took the

goat, which was the sin offering for the people, and slew it, and offered it for sin, as the first.

tauit hircum expiatoque altari 9:16 fecit holocaustum 9:17

9:16 And he brought the burnt offering, and offered it according to the manner. 9:17 And he brought the meat offering, and took an handful thereof, and burnt it upon the altar, beside the burnt sacrifice of the morning.

addens in sacrificio libamenta quae pariter offeruntur et adolens ea super altare absque caerimoniis holocausti matutini 9:18 immolauit et bouem atque arietem hostias pacifi- 9:18 He slew also the bullock and the ram for a sacrifice

of peace offerings, which was for the people: and Aaron’s sons presented unto him the blood, which he sprinkled

cas populi obtuleruntque ei filii sui sanguinem quem fudit upon the altar round about,

Biblia Sacra Vulgata

9:19 And the fat of the bullock and of the ram, the rump, super altare in circuitu 9:19 adipes autem bouis et caudam

and that which covereth the inwards, and the kidneys, and the caul above the liver:

arietis renunculosque cum adipibus suis et reticulum iecoris

9:20 And they put the fat upon the breasts, and he burnt 9:20 posuerunt super pectora cumque cremati essent adipes

the fat upon the altar:

9:21 And the breasts and the right shoulder Aaron waved in altari 9:21 pectora eorum et armos dextros separauit Aa-

for a wave offering before the LORD; as Moses commanded.

ron eleuans coram Domino sicut praeceperat Moses 9:22

9:22 And Aaron lifted up his hand toward the people, and blessed them, and came down from offering of the sin offering, and the burnt offering, and peace offerings.

et tendens manum contra populum benedixit eis sicque conpletis hostiis pro peccato et holocaustis et pacificis descendit

9:23 And Moses and Aaron went into the tabernacle of the 9:23 ingressi autem Moses et Aaron tabernaculum testimonii

congregation, and came out, and blessed the people: and the glory of the LORD appeared unto all the people.

et deinceps egressi benedixerunt populo apparuitque gloria

9:24 And there came a fire out from before the LORD, Domini omni multitudini 9:24 et ecce egressus ignis a Do-

and consumed upon the altar the burnt offering and the fat: which when all the people saw, they shouted, and fell

on their faces. mino deuorauit holocaustum et adipes qui erant super altare quod cum uidissent turbae laudauerunt Dominum ruentes in facies suas

10:1 And Nadab and Abihu, the sons of Aaron, took either 10:1 arreptisque Nadab et Abiu filii Aaron turibulis posue-

of them his censer, and put fire therein, and put incense thereon, and offered strange fire before the LORD, which

he commanded them not. runt ignem et incensum desuper offerentes coram Domino 10:2 And there went out fire from the LORD, and de- ignem alienum quod eis praeceptum non erat 10:2 egres-

voured them, and they died before the LORD.

susque ignis a Domino deuorauit eos et mortui sunt coram

10:3 Then Moses said unto Aaron, This is it that the Domino 10:3 dixitque Moses ad Aaron hoc est quod locu-

LORD spake, saying, I will be sanctified in them that come nigh me, and before all the people I will be glori-

fied. And Aaron held his peace. tus est Dominus sanctificabor in his qui adpropinquant mihi

et in conspectu omnis populi glorificabor quod audiens ta-

10:4 And Moses called Mishael and Elzaphan, the sons cuit Aaron 10:4 uocatis autem Moses Misahel et Elsaphan

of Uzziel the uncle of Aaron, and said unto them, Come near, carry your brethren from before the sanctuary out of

the camp. filios Ozihel patrui Aaron ait ad eos ite et colligite frat-

res uestros de conspectu sanctuarii et asportate extra castra

10:5 So they went near, and carried them in their coats out 10:5 confestimque pergentes tulerunt eos sicut iacebant ue-

of the camp; as Moses had said.

stitos lineis tunicis et eiecerunt foras ut sibi fuerat impera-

10:6 And Moses said unto Aaron, and unto Eleazar and tum 10:6 locutus est Moses ad Aaron et ad Eleazar atque

unto Ithamar, his sons, Uncover not your heads, neither rend your clothes; lest ye die, and lest wrath come upon all the people: but let your brethren, the whole house of Israel, bewail the burning which the LORD hath kindled.

Ithamar filios eius capita uestra nolite nudare et uestimenta nolite scindere ne forte moriamini et super omnem coetum oriatur indignatio fratres uestri et omnis domus Israhel plan-

10:7 And ye shall not go out from the door of the taber- gant incendium quod Dominus suscitauit 10:7 uos autem

nacle of the congregation, lest ye die: for the anointing oil of the LORD is upon you. And they did according to the

word of Moses. non egredimini fores tabernaculi alioquin peribitis oleum

quippe sanctae unctionis est super uos qui fecerunt omnia

10:8 And the LORD spake unto Aaron, saying, iuxta praeceptum Mosi 10:8 dixit quoque Dominus ad Aa-

Liber Leuiticus

ron 10:9 uinum et omne quod inebriare potest non bibetis tu 10:9 Do not drink wine nor strong drink, thou, nor thy sons

with thee, when ye go into the tabernacle of the congregation, lest ye die: it shall be a statute for ever throughout

et filii tui quando intratis tabernaculum testimonii ne moria- your generations: mini quia praeceptum est sempiternum in generationes uestras 10:10 et ut habeatis scientiam discernendi inter sanc- 10:10 And that ye may put difference between holy and

unholy, and between unclean and clean;

tum et profanum inter pollutum et mundum 10:11 doceatis- 10:11 And that ye may teach the children of Israel all the

statutes which the LORD hath spoken unto them by the hand of Moses.

que filios Israhel omnia legitima mea quae locutus est Dominus ad eos per manum Mosi 10:12 locutusque est Moses 10:12 And Moses spake unto Aaron, and unto Eleazar and

unto Ithamar, his sons that were left, Take the meat offering that remaineth of the offerings of the LORD made by fire, and eat it without leaven beside the altar: for it is most holy:

ad Aaron et ad Eleazar atque Ithamar filios eius qui residui erant tollite sacrificium quod remansit de oblatione Domini et comedite illud absque fermento iuxta altare quia sanctum sanctorum est 10:13 comedetis autem in loco sancto quod 10:13 And ye shall eat it in the holy place, because it is

thy due, and thy sons’ due, of the sacrifices of the LORD made by fire: for so I am commanded.

datum est tibi et filiis tuis de oblationibus Domini sicut praeceptum est mihi 10:14 pectusculum quoque quod oblatum 10:14 And the wave breast and heave shoulder shall ye

eat in a clean place; thou, and thy sons, and thy daughters with thee: for they be thy due, and thy sons’ due, which are given out of the sacrifices of peace offerings of the children of Israel.

est et armum qui separatus est edetis in loco mundissimo tu et filii tui ac filiae tuae tecum tibi enim ac liberis tuis reposita sunt de hostiis salutaribus filiorum Israhel 10:15 eo 10:15 The heave shoulder and the wave breast shall they

bring with the offerings made by fire of the fat, to wave it for a wave offering before the LORD; and it shall be thine, and thy sons’ with thee, by a statute for ever; as the LORD hath commanded.

quod armum et pectus et adipes qui cremantur in altari eleuauerint coram Domino et pertineant ad te et ad filios tuos lege perpetua sicut praecepit Dominus 10:16 inter haec hir- 10:16 And Moses diligently sought the goat of the sin of-

fering, and, behold, it was burnt: and he was angry with Eleazar and Ithamar, the sons of Aaron which were left

cum qui oblatus fuerat pro peccato cum quaereret Moses alive, saying, exustum repperit iratusque contra Eleazar et Ithamar filios Aaron qui remanserant ait 10:17 cur non comedistis hos- 10:17 Wherefore have ye not eaten the sin offering in the

holy place, seeing it is most holy, and God hath given it you to bear the iniquity of the congregation, to make

tiam pro peccato in loco sancto quae sancta sanctorum est et atonement for them before the LORD? data uobis ut portetis iniquitatem multitudinis et rogetis pro ea in conspectu Domini 10:18 praesertim cum de sanguine 10:18 Behold, the blood of it was not brought in within

the holy place: ye should indeed have eaten it in the holy place, as I commanded.

illius non sit inlatum intra sancta et comedere eam debueritis in sanctuario sicut praeceptum est mihi 10:19 respondit 10:19 And Aaron said unto Moses, Behold, this day have

they offered their sin offering and their burnt offering before the LORD; and such things have befallen me: and if I had eaten the sin offering to day, should it have been accepted in the sight of the LORD?

Aaron oblata est hodie uictima pro peccato et holocaustum coram Domino mihi autem accidit quod uides quomodo potui comedere eam aut placere Domino in caerimoniis mente lugubri 10:20 quod cum audisset Moses recepit satisfactio- 10:20 And when Moses heard that, he was content. nem

11:1 locutus est Dominus ad Mosen et Aaron dicens 11:2 11:1 And the LORD spake unto Moses and to Aaron, say-

ing unto them,

11:2 Speak unto the children of Israel, saying, These are the beasts which ye shall eat among all the beasts that are on the earth.

CLXXIX

Biblia Sacra Vulgata

dicite filiis Israhel haec sunt animalia quae comedere de-

11:3 Whatsoever parteth the hoof, and is clovenfooted, betis de cunctis animantibus terrae 11:3 omne quod habet

and cheweth the cud, among the beasts, that shall ye eat. 11:4 Nevertheless these shall ye not eat of them that chew diuisam ungulam et ruminat in pecoribus comedetis 11:4 the cud, or of them that divide the hoof: as the camel, because he cheweth the cud, but divideth not the hoof; he

is unclean unto you. quicquid autem ruminat quidem et habet ungulam sed non

diuidit eam sicut camelus et cetera non comedetis illud et

11:5 And the coney, because he cheweth the cud, but di- inter inmunda reputabitis 11:5 chyrogryllius qui ruminat

videth not the hoof; he is unclean unto you.

11:6 And the hare, because he cheweth the cud, but di- ungulamque non diuidit inmundus est 11:6 lepus quoque

videth not the hoof; he is unclean unto you.

11:7 And the swine, though he divide the hoof, and be nam et ipse ruminat sed ungulam non diuidit 11:7 et sus

clovenfooted, yet he cheweth not the cud; he is unclean to you.

qui cum ungulam diuidat non ruminat 11:8 horum carni-

11:8 Of their flesh shall ye not eat, and their carcase shall ye not touch; they are unclean to you.

bus non uescemini nec cadauera contingetis quia inmunda

11:9 These shall ye eat of all that are in the waters: what- sunt uobis 11:9 haec sunt quae gignuntur in aquis et uesci soever hath fins and scales in the waters, in the seas, and in the rivers, them shall ye eat.

licitum est omne quod habet pinnulas et squamas tam in

11:10 And all that have not fins and scales in the seas, mari quam in fluminibus et stagnis comedetis 11:10 quic-

and in the rivers, of all that move in the waters, and of any living thing which is in the waters, they shall be an

abomination unto you: quid autem pinnulas et squamas non habet eorum quae in 11:11 They shall be even an abomination unto you; ye aquis mouentur et uiuunt abominabile uobis 11:11 et exec-

shall not eat of their flesh, but ye shall have their carcases in abomination.

randum erit carnes eorum non comedetis et morticina uita-

11:12 Whatsoever hath no fins nor scales in the waters, bitis 11:12 cuncta quae non habent pinnulas et squamas in

that shall be an abomination unto you.

11:13 And these are they which ye shall have in abomina- aquis polluta erunt 11:13 haec sunt quae de auibus come-

tion among the fowls; they shall not be eaten, they are an abomination: the eagle, and the ossifrage, and the ospray,

dere non debetis et uitanda sunt uobis aquilam et grypem

11:14 And the vulture, and the kite after his kind; et alietum 11:14 miluum ac uulturem iuxta genus suum 11:15 Every raven after his kind; 11:15 et omne coruini generis in similitudinem suam 11:16

11:16 And the owl, and the night hawk, and the cuckow, and the hawk after his kind,

strutionem et noctuam et larum et accipitrem iuxta genus

11:17 And the little owl, and the cormorant, and the great suum 11:17 bubonem et mergulum et ibin 11:18 cycnum owl, 11:18 And the swan, and the pelican, and the gier eagle,

et onocrotalum et porphirionem 11:19 erodionem et chara-

11:19 And the stork, the heron after her kind, and the lapwing, and the bat.

drion iuxta genus suum opupam quoque et uespertilionem

11:20 All fowls that creep, going upon all four, shall be an 11:20 omne de uolucribus quod graditur super quattuor peabomination unto you.

11:21 Yet these may ye eat of every flying creeping thing des abominabile erit uobis 11:21 quicquid autem ambulat

that goeth upon all four, which have legs above their feet, to leap withal upon the earth;

quidem super quattuor pedes sed habet longiora retro crura

11:22 Even these of them ye may eat; the locust after his per quae salit super terram 11:22 comedere debetis ut est

kind, and the bald locust after his kind, and the beetle after his kind, and the grasshopper after his kind.

brucus in genere suo et attacus atque ophiomachus ac lu-

11:23 But all other flying creeping things, which have four custa singula iuxta genus suum 11:23 quicquid autem ex feet, shall be an abomination unto you.

uolucribus quattuor tantum habet pedes execrabile erit uobis

11:24 And for these ye shall be unclean: whosoever 11:24 et quicumque morticina eorum tetigerit polluetur et erit

toucheth the carcase of them shall be unclean until the even.

Liber Leuiticus

inmundus usque ad uesperum 11:25 et si necesse fuerit ut 11:25 And whosoever beareth ought of the carcase of them

shall wash his clothes, and be unclean until the even.

portet quippiam horum mortuum lauabit uestimenta sua et

inmundus erit usque ad solis occasum 11:26 omne animal 11:26 The carcases of every beast which divideth the hoof, and is not clovenfooted, nor cheweth the cud, are unclean unto you: every one that toucheth them shall be unclean.

quod habet quidem ungulam sed non diuidit eam nec ruminat inmundum erit et quicquid tetigerit illud contaminabitur 11:27 quod ambulat super manus ex cunctis animantibus quae 11:27 And whatsoever goeth upon his paws, among all

manner of beasts that go on all four, those are unclean unto you: whoso toucheth their carcase shall be unclean

incedunt quadrupedia inmundum erit qui tetigerit morticina until the even. eorum polluetur usque ad uesperum 11:28 et qui portauerit 11:28 And he that beareth the carcase of them shall wash

his clothes, and be unclean until the even: they are unclean unto you.

huiuscemodi cadauera lauabit uestimenta sua et inmundus erit usque ad uesperum quia omnia haec inmunda sunt uobis 11:29 hoc quoque inter polluta reputabitur de his quae mo- 11:29 These also shall be unclean unto you among the

creeping things that creep upon the earth; the weasel, and the mouse, and the tortoise after his kind,

uentur in terra mustela et mus et corcodillus singula iuxta genus suum 11:30 migale et cameleon et stelio ac lacerta 11:30 And the ferret, and the chameleon, and the lizard,

and the snail, and the mole.

et talpa 11:31 omnia haec inmunda sunt qui tetigerit morti- 11:31 These are unclean to you among all that creep:

whosoever doth touch them, when they be dead, shall be unclean until the even.

cina eorum inmundus erit usque ad uesperum 11:32 et su-

11:32 And upon whatsoever any of them, when they are dead, doth fall, it shall be unclean; whether it be any vessel of wood, or raiment, or skin, or sack, whatsoever vessel it be, wherein any work is done, it must be put into water, and it shall be unclean until the even; so it shall be cleansed.

per quod ceciderit quicquam de morticinis eorum polluetur tam uas ligneum et uestimentum quam pelles et cilicia et in quocumque fit opus tinguentur aqua et polluta erunt usque ad uesperum et sic postea mundabuntur 11:33 uas au- 11:33 And every earthen vessel, whereinto any of them

falleth, whatsoever is in it shall be unclean; and ye shall break it.

tem fictile in quo horum quicquam intro ceciderit polluetur et idcirco frangendum est 11:34 omnis cibus quem come- 11:34 Of all meat which may be eaten, that on which such

water cometh shall be unclean: and all drink that may be drunk in every such vessel shall be unclean.

ditis si fusa fuerit super eum aqua inmundus erit et omne liquens quod bibitur de uniuerso uase inmundum erit 11:35 11:35 And every thing whereupon any part of their carcase

falleth shall be unclean; whether it be oven, or ranges for pots, they shall be broken down: for they are unclean and

et quicquid de morticinis istiusmodi ceciderit super illud in- shall be unclean unto you. mundum erit siue clibani siue cytropodes destruentur et inmundi erunt 11:36 fontes uero et cisternae et omnis aqua- 11:36 Nevertheless a fountain or pit, wherein there is

plenty of water, shall be clean: but that which toucheth their carcase shall be unclean.

rum congregatio munda erit qui morticinum eorum tetigerit polluetur 11:37 si ceciderint super sementem non polluent 11:37 And if any part of their carcase fall upon any sowing

seed which is to be sown, it shall be clean.

eam 11:38 sin autem quispiam aqua sementem perfuderit et 11:38 But if any water be put upon the seed, and any part

of their carcase fall thereon, it shall be unclean unto you.

postea morticinis tacta fuerit ilico polluetur 11:39 si mor- 11:39 And if any beast, of which ye may eat, die; he

that toucheth the carcase thereof shall be unclean until the even.

tuum fuerit animal quod licet uobis comedere qui cadauer eius tetigerit inmundus erit usque ad uesperum 11:40 et qui 11:40 And he that eateth of the carcase of it shall wash his

clothes, and be unclean until the even: he also that beareth the carcase of it shall wash his clothes, and be unclean

comederit ex eo quippiam siue portauerit lauabit uestimenta until the even.

Biblia Sacra Vulgata

11:41 And every creeping thing that creepeth upon the sua et inmundus erit usque ad uesperum 11:41 omne quod earth shall be an abomination; it shall not be eaten.

reptat super terram abominabile erit nec adsumetur in cibum

11:42 Whatsoever goeth upon the belly, and whatsoever 11:42 quicquid super pectus quadrupes graditur et multos ha-

goeth upon all four, or whatsoever hath more feet among all creeping things that creep upon the earth, them ye shall

not eat; for they are an abomination. bet pedes siue per humum trahitur non comedetis quia abo-11:43 Ye shall not make yourselves abominable with minabile est 11:43 nolite contaminare animas uestras nec

any creeping thing that creepeth, neither shall ye make yourselves unclean with them, that ye should be defiled

thereby. tangatis quicquam eorum ne inmundi sitis 11:44 ego enim

11:44 For I am the LORD your God: ye shall therefore sanctify yourselves, and ye shall be holy; for I am holy: neither shall ye defile yourselves with any manner of creeping thing that creepeth upon the earth.

sum Dominus Deus uester sancti estote quoniam et ego sanctus sum ne polluatis animas uestras in omni reptili quod mo-

11:45 For I am the LORD that bringeth you up out of the uetur super terram 11:45 ego sum Dominus qui eduxi uos

land of Egypt, to be your God: ye shall therefore be holy, for I am holy.

de terra Aegypti ut essem uobis in Deum sancti eritis quia

11:46 This is the law of the beasts, and of the fowl, and et ego sanctus sum 11:46 ista est lex animantium et uoluc-

of every living creature that moveth in the waters, and of every creature that creepeth upon the earth:

rum et omnis animae uiuentis quae mouetur in aqua et reptat

11:47 To make a difference between the unclean and the in terra 11:47 ut differentias noueritis mundi et inmundi et

clean, and between the beast that may be eaten and the beast that may not be eaten.

sciatis quid comedere et quid respuere debeatis

12:1 And the LORD spake unto Moses, saying, 12:1 locutus est Dominus ad Mosen dicens 12:2 loquere

12:2 Speak unto the children of Israel, saying, If a woman have conceived seed, and born a man child: then she shall be unclean seven days; according to the days of the separation for her infirmity shall she be unclean.

filiis Israhel et dices ad eos mulier si suscepto semine pepererit masculum inmunda erit septem diebus iuxta dies sepa-

12:3 And in the eighth day the flesh of his foreskin shall rationis menstruae 12:3 et die octauo circumcidetur infan-

be circumcised.

12:4 And she shall then continue in the blood of her pu- tulus 12:4 ipsa uero triginta tribus diebus manebit in san-

rifying three and thirty days; she shall touch no hallowed thing, nor come into the sanctuary, until the days of her

purifying be fulfilled. guine purificationis suae omne sanctum non tanget nec in-

gredietur sanctuarium donec impleantur dies purificationis

12:5 But if she bear a maid child, then she shall be unclean eius 12:5 sin autem feminam pepererit inmunda erit duatwo weeks, as in her separation: and she shall continue in the blood of her purifying threescore and six days.

bus ebdomadibus iuxta ritum fluxus menstrui et sexaginta

12:6 And when the days of her purifying are fulfilled, for ac sex diebus manebit in sanguine purificationis suae 12:6

a son, or for a daughter, she shall bring a lamb of the first year for a burnt offering, and a young pigeon, or a turtledove, for a sin offering, unto the door of the tabernacle of the congregation, unto the priest:

cumque expleti fuerint dies purificationis eius pro filio siue pro filia deferet agnum anniculum in holocaustum et pullum columbae siue turturem pro peccato ad ostium taberna-

12:7 Who shall offer it before the LORD, and make an culi testimonii et tradet sacerdoti 12:7 qui offeret illa coram

atonement for her; and she shall be cleansed from the issue of her blood. This is the law for her that hath born a male

or a female. Domino et rogabit pro ea et sic mundabitur a profluuio san-12:8 And if she be not able to bring a lamb, then she shall guinis sui ista est lex parientis masculum ac feminam 12:8

bring two turtles, or two young pigeons; the one for the burnt offering, and the other for a sin offering: and the priest shall make an atonement for her, and she shall be clean.

quod si non inuenerit manus eius nec potuerit offerre agnum sumet duos turtures uel duos pullos columbae unum in holocaustum et alterum pro peccato orabitque pro ea sacerdos

Liber Leuiticus

et sic mundabitur

13:1 locutus est Dominus ad Mosen et Aaron dicens 13:2 13:1 And the LORD spake unto Moses and Aaron, saying,

13:2 When a man shall have in the skin of his flesh a rising, a scab, or bright spot, and it be in the skin of his flesh like the plague of leprosy; then he shall be brought unto Aaron the priest, or unto one of his sons the priests:

homo in cuius carne et cute ortus fuerit diuersus color siue pustula aut quasi lucens quippiam id est plaga leprae adducetur ad Aaron sacerdotem uel ad unum quemlibet filiorum eius 13:3 qui cum uiderit lepram in cute et pilos in album 13:3 And the priest shall look on the plague in the skin of

the flesh: and when the hair in the plague is turned white, and the plague in sight be deeper than the skin of his flesh, it is a plague of leprosy: and the priest shall look on him, and pronounce him unclean.

mutatos colorem ipsamque speciem leprae humiliorem cute et carne reliqua plaga leprae est et ad arbitrium eius separabitur 13:4 sin autem lucens candor fuerit in cute nec hu- 13:4 If the bright spot be white in the skin of his flesh, and

in sight be not deeper than the skin, and the hair thereof be not turned white; then the priest shall shut up him that

milior carne reliqua et pili coloris pristini recludet eum sa- hath the plague seven days: cerdos septem diebus 13:5 et considerabit die septimo et si- 13:5 And the priest shall look on him the seventh day: and,

behold, if the plague in his sight be at a stay, and the plague spread not in the skin; then the priest shall shut him up

quidem lepra ultra non creuerit nec transierit in cute priores seven days more: terminos rursum includet eum septem diebus aliis 13:6 et 13:6 And the priest shall look on him again the seventh

day: and, behold, if the plague be somewhat dark, and the plague spread not in the skin, the priest shall pronounce him clean: it is but a scab: and he shall wash his clothes, and be clean.

die septimo contemplabitur si obscurior fuerit lepra et non creuerit in cute mundabit eum quia scabies est lauabitque homo uestimenta sua et mundus erit 13:7 quod si postquam 13:7 But if the scab spread much abroad in the skin, after

that he hath been seen of the priest for his cleansing, he shall be seen of the priest again.

a sacerdote uisus est et redditus munditiae iterum lepra creuerit adducetur ad eum 13:8 et inmunditiae condemnabitur 13:8 And if the priest see that, behold, the scab spreadeth

in the skin, then the priest shall pronounce him unclean: it is a leprosy.

13:9 plaga leprae si fuerit in homine adducetur ad sacerdo-

13:9 When the plague of leprosy is in a man, then he shall be brought unto the priest;

tem 13:10 et uidebit eum cumque color albus in cute fuerit 13:10 And the priest shall see him: and, behold, if the rising be white in the skin, and it have turned the hair white, and there be quick raw flesh in the rising;

et capillorum mutarit aspectum ipsa quoque caro uiua apparuerit 13:11 lepra uetustissima iudicabitur atque inolita cuti 13:11 It is an old leprosy in the skin of his flesh, and the

priest shall pronounce him unclean, and shall not shut him up: for he is unclean.

contaminabit itaque eum sacerdos et non recludet quia perspicue inmunditia est 13:12 sin autem effloruerit discurrens 13:12 And if a leprosy break out abroad in the skin, and the

leprosy cover all the skin of him that hath the plague from his head even to his foot, wheresoever the priest looketh;

lepra in cute et operuerit omnem carnem a capite usque ad pedes quicquid sub aspectu oculorum cadit 13:13 conside- 13:13 Then the priest shall consider: and, behold, if the

leprosy have covered all his flesh, he shall pronounce him clean that hath the plague: it is all turned white: he is

rabit eum sacerdos et teneri lepra mundissima iudicabit eo clean. quod omnis in candorem uersa sit et idcirco homo mundus erit 13:14 quando uero caro uiuens in eo apparuerit 13:15 13:14 But when raw flesh appeareth in him, he shall be

unclean.

13:15 And the priest shall see the raw flesh, and pronounce him to be unclean: for the raw flesh is unclean: it is a leprosy.

tunc sacerdotis iudicio polluetur et inter inmundos reputabitur caro enim uiua si lepra aspergatur inmunda est 13:16 13:16 Or if the raw flesh turn again, and be changed unto

white, he shall come unto the priest;

quod si rursum uersa fuerit in alborem et totum hominem operuerit 13:17 considerabit eum sacerdos et mundum esse 13:17 And the priest shall see him: and, behold, if the

plague be turned into white; then the priest shall pronounce him clean that hath the plague: he is clean.

Biblia Sacra Vulgata

13:18 The flesh also, in which, even in the skin thereof, decernet 13:18 caro et cutis in qua ulcus natum est et sa-

was a boil, and is healed,

13:19 And in the place of the boil there be a white rising, natum 13:19 et in loco ulceris cicatrix apparuerit alba siue

or a bright spot, white, and somewhat reddish, and it be shewed to the priest;

subrufa adducetur homo ad sacerdotem 13:20 qui cum ui-

13:20 And if, when the priest seeth it, behold, it be in sight lower than the skin, and the hair thereof be turned white; the priest shall pronounce him unclean: it is a plague of leprosy broken out of the boil.

derit locum leprae humiliorem carne reliqua et pilos uersos in candorem contaminabit eum plaga enim leprae orta est in

13:21 But if the priest look on it, and, behold, there be no ulcere 13:21 quod si pilus coloris est pristini et cicatrix su-

white hairs therein, and if it be not lower than the skin, but be somewhat dark; then the priest shall shut him up seven

days: bobscura et uicina carne non est humilior recludet eum sep-13:22 And if it spread much abroad in the skin, then the tem diebus 13:22 et siquidem creuerit adiudicabit eum lep-

priest shall pronounce him unclean: it is a plague.

13:23 But if the bright spot stay in his place, and spread rae 13:23 sin autem steterit in loco suo ulceris est cicatrix et

not, it is a burning boil; and the priest shall pronounce him clean.

homo mundus erit 13:24 caro et cutis quam ignis exuserit et

13:24 Or if there be any flesh, in the skin whereof there is a hot burning, and the quick flesh that burneth have a white bright spot, somewhat reddish, or white;

sanata albam siue rufam habuerit cicatricem 13:25 conside-

13:25 Then the priest shall look upon it: and, behold, if the hair in the bright spot be turned white, and it be in sight deeper than the skin; it is a leprosy broken out of the burning: wherefore the priest shall pronounce him unclean: it is the plague of leprosy.

rabit eam sacerdos et ecce uersa est in alborem et locus eius reliqua cute humilior contaminabit eum quia plaga leprae in

13:26 But if the priest look on it, and, behold, there be no cicatrice orta est 13:26 quod si pilorum color non fuerit in-

white hair in the bright spot, and it be no lower than the other skin, but be somewhat dark; then the priest shall shut

him up seven days: mutatus nec humilior plaga carne reliqua et ipsa leprae spe-13:27 And the priest shall look upon him the seventh day: cies fuerit subobscura recludet eum septem diebus 13:27 et

and if it be spread much abroad in the skin, then the priest shall pronounce him unclean: it is the plague of leprosy.

die septimo contemplabitur si creuerit in cute lepra contami-

13:28 And if the bright spot stay in his place, and spread nabit eum 13:28 sin autem in loco suo candor steterit non not in the skin, but it be somewhat dark; it is a rising of the burning, and the priest shall pronounce him clean: for

it is an inflammation of the burning. satis clarus plaga conbustionis est et idcirco mundabitur quia 13:29 If a man or woman have a plague upon the head or cicatrix est conbusturae 13:29 uir siue mulier in cuius capite

the beard;

13:30 Then the priest shall see the plague: and, behold, if uel barba germinarit lepra uidebit eos sacerdos 13:30 et si-

it be in sight deeper than the skin; and there be in it a yellow thin hair; then the priest shall pronounce him unclean:

it is a dry scall, even a leprosy upon the head or beard. quidem humilior fuerit locus carne reliqua et capillus flauus

solitoque subtilior contaminabit eos quia lepra capitis ac bar-

13:31 And if the priest look on the plague of the scall, bae est 13:31 sin autem uiderit et locum maculae aequalem

and, behold, it be not in sight deeper than the skin, and that there is no black hair in it; then the priest shall shut up

him that hath the plague of the scall seven days: uicinae carni et capillum nigrum recludet eos septem diebus 13:32 And in the seventh day the priest shall look on the 13:32 et die septimo intuebitur si non creuerit macula et ca-

plague: and, behold, if the scall spread not, and there be in it no yellow hair, and the scall be not in sight deeper than

the skin; pillus sui coloris est et locus plagae carni reliquae aequa-

13:33 He shall be shaven, but the scall shall he not shave; lis 13:33 radetur homo absque loco maculae et includetur and the priest shall shut up him that hath the scall seven days more:

septem diebus aliis 13:34 si die septimo uisa fuerit stetisse

13:34 And in the seventh day the priest shall look on the scall: and, behold, if the scall be not spread in the skin, nor be in sight deeper than the skin; then the priest shall pronounce him clean: and he shall wash his clothes, and be clean.

plaga in loco suo nec humilior carne reliqua mundabit eum

13:35 But if the scall spread much in the skin after his lotisque uestibus mundus erit 13:35 sin autem post emun-

cleansing;

13:36 Then the priest shall look on him: and, behold, if dationem rursus creuerit macula in cute 13:36 non quae-

the scall be spread in the skin, the priest shall not seek for yellow hair; he is unclean.

ret amplius utrum capillus in flauum colorem sit commu-

Liber Leuiticus

tatus quia aperte inmundus est 13:37 porro si steterit ma- 13:37 But if the scall be in his sight at a stay, and that

there is black hair grown up therein; the scall is healed, he is clean: and the priest shall pronounce him clean.

cula et capilli nigri fuerint nouerit hominem esse sanatum et confidenter eum pronuntiet mundum 13:38 uir et mulier in 13:38 If a man also or a woman have in the skin of their

flesh bright spots, even white bright spots;

cuius cute candor apparuerit 13:39 intuebitur eos sacerdos 13:39 Then the priest shall look: and, behold, if the bright

spots in the skin of their flesh be darkish white; it is a freckled spot that groweth in the skin; he is clean.

si deprehenderit subobscurum alborem lucere in cute sciat non esse lepram sed maculam coloris candidi et hominem mundum 13:40 uir de cuius capite capilli fluunt caluus ac 13:40 And the man whose hair is fallen off his head, he is

bald; yet is he clean.

mundus est 13:41 et si a fronte ceciderint pili recaluaster et 13:41 And he that hath his hair fallen off from the part of

his head toward his face, he is forehead bald: yet is he clean.

mundus est 13:42 sin autem in caluitio siue in recaluatione

13:42 And if there be in the bald head, or bald forehead, a white reddish sore; it is a leprosy sprung up in his bald head, or his bald forehead.

albus uel rufus color fuerit exortus 13:43 et hoc sacerdos

13:43 Then the priest shall look upon it: and, behold, if the rising of the sore be white reddish in his bald head, or in his bald forehead, as the leprosy appeareth in the skin of the flesh;

uiderit condemnabit eum haut dubiae leprae quae orta est in caluitio 13:44 quicumque ergo maculatus fuerit lepra et 13:44 He is a leprous man, he is unclean: the priest shall

pronounce him utterly unclean; his plague is in his head.

separatus ad arbitrium sacerdotis 13:45 habebit uestimenta 13:45 And the leper in whom the plague is, his clothes

shall be rent, and his head bare, and he shall put a covering upon his upper lip, and shall cry, Unclean, unclean.

dissuta caput nudum os ueste contectum contaminatum ac sordidum se clamabit 13:46 omni tempore quo leprosus est 13:46 All the days wherein the plague shall be in him he

shall be defiled; he is unclean: he shall dwell alone; without the camp shall his habitation be.

et inmundus solus habitabit extra castra 13:47 uestis lanea

13:47 The garment also that the plague of leprosy is in, whether it be a woollen garment, or a linen garment;

siue linea quae lepram habuerit 13:48 in stamine atque sub- _{13:48 Whether it be in the warp, or woof; of linen, or of}

woollen; whether in a skin, or in any thing made of skin;

temine aut certe pellis uel quicquid ex pelle confectum est 13:49 si alba aut rufa macula fuerit infecta lepra reputabi- 13:49 And if the plague be greenish or reddish in the gar-

ment, or in the skin, either in the warp, or in the woof, or in any thing of skin; it is a plague of leprosy, and shall be

tur ostendeturque sacerdoti 13:50 qui consideratam reclu- shewed unto the priest:

13:50 And the priest shall look upon the plague, and shut up it that hath the plague seven days:

det septem diebus 13:51 et die septimo rursus aspiciens si

13:51 And he shall look on the plague on the seventh day: if the plague be spread in the garment, either in the warp, or in the woof, or in a skin, or in any work that is made of skin; the plague is a fretting leprosy; it is unclean.

creuisse deprehenderit lepra perseuerans est pollutum iudi-

cabit uestimentum et omne in quo fuerit inuenta 13:52 et 13:52 He shall therefore burn that garment, whether warp or woof, in woollen or in linen, or any thing of skin, wherein the plague is: for it is a fretting leprosy; it shall

idcirco conburetur flammis 13:53 quod si eam uiderit non be burnt in the fire.

13:53 And if the priest shall look, and, behold, the plague be not spread in the garment, either in the warp, or in the woof, or in any thing of skin;

creuisse 13:54 praecipiet et lauabunt id in quo lepra est re-

13:54 Then the priest shall command that they wash the thing wherein the plague is, and he shall shut it up seven days more:

cludetque illud septem diebus aliis 13:55 et cum uiderit fa-

13:55 And the priest shall look on the plague, after that it is washed: and, behold, if the plague have not changed his colour, and the plague be not spread; it is unclean; thou shalt burn it in the fire; it is fret inward, whether it be bare within or without.

ciem quidem pristinam non reuersam nec tamen creuisse lepram inmundum iudicabit et igne conburet eo quod infusa sit in superficie uestimenti uel per totum lepra 13:56 sin autem 13:56 And if the priest look, and, behold, the plague be

somewhat dark after the washing of it; then he shall rend it out of the garment, or out of the skin, or out of the warp,

obscurior fuerit locus leprae postquam uestis est lota abrum- or out of the woof: pet eum et a solido diuidet 13:57 quod si ultra apparuerit in 13:57 And if it appear still in the garment, either in the

warp, or in the woof, or in any thing of skin; it is a spreading plague: thou shalt burn that wherein the plague is with

his locis quae prius inmaculata erant lepra uolatilis et uaga fire. debet igne conburi 13:58 si cessauerit lauabit ea quae pura 13:58 And the garment, either warp, or woof, or what-

soever thing of skin it be, which thou shalt wash, if the plague be departed from them, then it shall be washed the second time, and shall be clean.

Biblia Sacra Vulgata

13:59 This is the law of the plague of leprosy in a garment sunt secundo et munda erunt 13:59 ista est lex leprae ueof woollen or linen, either in the warp, or woof, or any thing of skins, to pronounce it clean, or to pronounce it

unclean. stimenti lanei et linei staminis atque subteminis omnisque supellectilis pelliciae quomodo mundari debeat uel contaminari

14:1 And the LORD spake unto Moses, saying, 14:1 locutusque est Dominus ad Mosen dicens 14:2 hic

14:2 This shall be the law of the leper in the day of his cleansing: He shall be brought unto the priest:

est ritus leprosi quando mundandus est adducetur ad sacer-

14:3 And the priest shall go forth out of the camp; and the dotem 14:3 qui egressus e castris cum inuenerit lepram esse

priest shall look, and, behold, if the plague of leprosy be healed in the leper;

mundatam 14:4 praecipiet ei qui purificatur ut offerat pro

14:4 Then shall the priest command to take for him that is to be cleansed two birds alive and clean, and cedar wood, and scarlet, and hyssop:

se duos passeres uiuos quos uesci licitum est et lignum ce-

14:5 And the priest shall command that one of the birds be drinum uermiculumque et hysopum 14:5 et unum e passekilled in an earthen vessel over running water:

ribus immolari iubebit in uase fictili super aquas uiuentes

14:6 As for the living bird, he shall take it, and the cedar 14:6 alium autem uiuum cum ligno cedrino et cocco et hywood, and the scarlet, and the hyssop, and shall dip them and the living bird in the blood of the bird that was killed

over the running water: sopo tinguet in sanguine passeris immolati 14:7 quo asper-

14:7 And he shall sprinkle upon him that is to be cleansed from the leprosy seven times, and shall pronounce him clean, and shall let the living bird loose into the open field.

get illum qui mundandus est septies ut iure purgetur et dimit-

14:8 And he that is to be cleansed shall wash his clothes, tet passerem uiuum ut in agrum auolet 14:8 cumque lauerit

and shave off all his hair, and wash himself in water, that he may be clean: and after that he shall come into the

camp, and shall tarry abroad out of his tent seven days. homo uestimenta sua radet omnes pilos corporis et lauabitur

aqua purificatusque ingredietur castra ita dumtaxat ut ma-

14:9 But it shall be on the seventh day, that he shall shave neat extra tabernaculum suum septem diebus 14:9 et die se-

all his hair off his head and his beard and his eyebrows, even all his hair he shall shave off: and he shall wash his clothes, also he shall wash his flesh in water, and he shall be clean.

ptimo radat capillos capitis barbamque et supercilia ac totius

14:10 And on the eighth day he shall take two he lambs corporis pilos et lotis rursum uestibus et corpore 14:10 die

without blemish, and one ewe lamb of the first year without blemish, and three tenth deals of fine flour for a meat

offering, mingled with oil, and one log of oil. octauo adsumet duos agnos inmaculatos et ouem anniculam

absque macula et tres decimas similae in sacrificium quae

14:11 And the priest that maketh him clean shall present conspersa sit oleo et seorsum olei sextarium 14:11 cumque

the man that is to be made clean, and those things, before the LORD, at the door of the tabernacle of the congrega-

tion: sacerdos purificans hominem statuerit eum et haec omnia 14:12 And the priest shall take one he lamb, and offer him coram Domino in ostio tabernaculi testimonii 14:12 tollet

for a trespass offering, and the log of oil, and wave them for a wave offering before the LORD:

agnum et offeret eum pro delicto oleique sextarium et obla-

14:13 And he shall slay the lamb in the place where he tis ante Dominum omnibus 14:13 immolabit agnum ubi im-

shall kill the sin offering and the burnt offering, in the holy place: for as the sin offering is the priest’s, so is the tres-

pass offering: it is most holy: molari solet hostia pro peccato et holocaustum id est in loco

sancto sicut enim pro peccato ita et pro delicto ad sacerdo-

14:14 And the priest shall take some of the blood of the tem pertinet hostia sancta sanctorum est 14:14 adsumens-

trespass offering, and the priest shall put it upon the tip of the right ear of him that is to be cleansed, and upon the thumb of his right hand, and upon the great toe of his right foot:

que sacerdos de sanguine hostiae quae immolata est pro delicto ponet super extremum auriculae dextrae eius qui mun-

14:15 And the priest shall take some of the log of oil, and datur et super pollices manus dextrae et pedis 14:15 et de

pour it into the palm of his own left hand:

Liber Leuiticus

olei sextario mittet in manum suam sinistram 14:16 tinguet- 14:16 And the priest shall dip his right finger in the oil

that is in his left hand, and shall sprinkle of the oil with his finger seven times before the LORD:

que digitum dextrum in eo et asperget septies contra Do-

minum 14:17 quod autem reliquum est olei in leua manu 14:17 And of the rest of the oil that is in his hand shall the priest put upon the tip of the right ear of him that is to be cleansed, and upon the thumb of his right hand, and upon the great toe of his right foot, upon the blood of the trespass offering:

fundet super extremum auriculae dextrae eius qui mundatur et super pollices manus ac pedis dextri et super sanguinem

qui fusus est pro delicto 14:18 et super caput eius 14:19 14:18 And the remnant of the oil that is in the priest’s hand he shall pour upon the head of him that is to be cleansed: and the priest shall make an atonement for him before the LORD.

14:19 And the priest shall offer the sin offering, and make an atonement for him that is to be cleansed from his uncleanness; and afterward he shall kill the burnt offering:

rogabitque pro eo coram Domino et faciet sacrificium pro peccato tunc immolabit holocaustum 14:20 et ponet illud

14:20 And the priest shall offer the burnt offering and the meat offering upon the altar: and the priest shall make an atonement for him, and he shall be clean.

in altari cum libamentis suis et homo rite mundabitur 14:21

14:21 And if he be poor, and cannot get so much; then he shall take one lamb for a trespass offering to be waved, to make an atonement for him, and one tenth deal of fine flour mingled with oil for a meat offering, and a log of oil;

quod si pauper est et non potest manus eius inuenire quae dicta sunt adsumet agnum pro delicto ad oblationem ut roget pro eo sacerdos decimamque partem similae conspersae oleo in sacrificium et olei sextarium 14:22 duosque turtures 14:22 And two turtledoves, or two young pigeons, such as

he is able to get; and the one shall be a sin offering, and the other a burnt offering.

siue duos pullos columbae quorum sit unus pro peccato et alter in holocaustum 14:23 offeretque ea die octauo purifi- 14:23 And he shall bring them on the eighth day for his

cleansing unto the priest, unto the door of the tabernacle of the congregation, before the LORD.

cationis suae sacerdoti ad ostium tabernaculi testimonii coram Domino 14:24 qui suscipiens agnum pro delicto et sex- 14:24 And the priest shall take the lamb of the trespass

offering, and the log of oil, and the priest shall wave them for a wave offering before the LORD:

tarium olei leuabit simul 14:25 immolatoque agno de san-

14:25 And he shall kill the lamb of the trespass offering, and the priest shall take some of the blood of the trespass offering, and put it upon the tip of the right ear of him that is to be cleansed, and upon the thumb of his right hand, and upon the great toe of his right foot:

guine eius ponet super extremum auriculae dextrae illius qui mundatur et super pollices manus eius ac pedis dextri 14:26 _{14:26 And the priest shall pour of the oil into the palm of}

his own left hand:

olei uero partem mittet in manum suam sinistram 14:27 in 14:27 And the priest shall sprinkle with his right finger

some of the oil that is in his left hand seven times before the LORD:

quo tinguens digitum dextrae manus asperget septies contra Dominum 14:28 tangetque extremum dextrae auriculae il- 14:28 And the priest shall put of the oil that is in his hand

upon the tip of the right ear of him that is to be cleansed, and upon the thumb of his right hand, and upon the great toe of his right foot, upon the place of the blood of the trespass offering:

lius qui mundatur et pollices manus ac pedis dextri in loco sanguinis qui effusus est pro delicto 14:29 reliquam autem 14:29 And the rest of the oil that is in the priest’s hand he

shall put upon the head of him that is to be cleansed, to make an atonement for him before the LORD.

partem olei quae est in sinistra manu mittet super caput purificati ut placet pro eo Dominum 14:30 et turturem siue pul- 14:30 And he shall offer the one of the turtledoves, or of

the young pigeons, such as he can get;

lum columbae offeret 14:31 unum pro delicto et alterum in 14:31 Even such as he is able to get, the one for a sin

offering, and the other for a burnt offering, with the meat offering: and the priest shall make an atonement for him

holocaustum cum libamentis suis 14:32 hoc est sacrificium that is to be cleansed before the LORD.

14:32 This is the law of him in whom is the plague of leprosy, whose hand is not able to get that which pertaineth to his cleansing.

leprosi qui habere non potest omnia in emundationem sui 14:33 locutus est Dominus ad Mosen et Aaron dicens 14:34 14:33 And the LORD spake unto Moses and unto Aaron,

saying,

14:34 When ye be come into the land of Canaan, which I give to you for a possession, and I put the plague of leprosy in a house of the land of your possession;

cum ingressi fueritis terram Chanaan quam ego dabo uobis in possessionem si fuerit plaga leprae in aedibus 14:35 ibit _{14:35 And he that owneth the house shall come and tell}

the priest, saying, It seemeth to me there is as it were a plague in the house:

Biblia Sacra Vulgata

cuius est domus nuntians sacerdoti et dicet quasi plaga lep-

14:36 Then the priest shall command that they empty the rae uidetur mihi esse in domo mea 14:36 at ille praecipiet

house, before the priest go into it to see the plague, that all that is in the house be not made unclean: and afterward

the priest shall go in to see the house: ut efferant uniuersa de domo priusquam ingrediatur eam et

uideat utrum lepra sit ne inmunda fiant omnia quae in domo

14:37 And he shall look on the plague, and, behold, if the sunt intrabitque postea ut consideret domus lepram 14:37 et

plague be in the walls of the house with hollow strakes, greenish or reddish, which in sight are lower than the wall;

cum uiderit in parietibus illius quasi ualliculas pallore siue

14:38 Then the priest shall go out of the house to the door rubore deformes et humiliores superficie reliqua 14:38 eg-

of the house, and shut up the house seven days:

redietur ostium domus et statim claudet eam septem diebus

14:39 And the priest shall come again the seventh day, and 14:39 reuersusque die septimo considerabit eam si inuenerit

shall look: and, behold, if the plague be spread in the walls of the house;

creuisse lepram 14:40 iubebit erui lapides in quibus lepra

14:40 Then the priest shall command that they take away the stones in which the plague is, and they shall cast them into an unclean place without the city:

est et proici eos extra ciuitatem in loco inmundo 14:41 do-

14:41 And he shall cause the house to be scraped within round about, and they shall pour out the dust that they

scrape off without the city into an unclean place: mum autem ipsam radi intrinsecus per circuitum et spargi

14:42 And they shall take other stones, and put them in the puluerem rasurae extra urbem in loco inmundo 14:42 laplace of those stones; and he shall take other morter, and shall plaister the house.

pidesque alios reponi pro his qui ablati fuerint et luto alio

14:43 And if the plague come again, and break out in the liniri domum 14:43 sin autem postquam eruti sunt lapides

house, after that he hath taken away the stones, and after he hath scraped the house, and after it is plaistered;

et puluis elatus et alia terra lita 14:44 ingressus sacerdos

14:44 Then the priest shall come and look, and, behold, if the plague be spread in the house, it is a fretting leprosy in the house; it is unclean.

uiderit reuersam lepram et parietes aspersos maculis lepra

14:45 And he shall break down the house, the stones of est perseuerans et inmunda domus 14:45 quam statim de-

it, and the timber thereof, and all the morter of the house; and he shall carry them forth out of the city into an unclean

place. struent et lapides eius ac ligna atque uniuersum puluerem 14:46 Moreover he that goeth into the house all the while proicient extra oppidum in loco inmundo 14:46 qui intra-

that it is shut up shall be unclean until the even.

uerit domum quando clausa est inmundus erit usque ad ue-

14:47 And he that lieth in the house shall wash his clothes; sperum 14:47 et qui dormierit in ea et comederit quippiam and he that eateth in the house shall wash his clothes.

14:48 And if the priest shall come in, and look upon it, lauabit uestimenta sua 14:48 quod si introiens sacerdos ui-

and, behold, the plague hath not spread in the house, after the house was plaistered: then the priest shall pronounce

the house clean, because the plague is healed. derit lepram non creuisse in domo postquam denuo lita est 14:49 And he shall take to cleanse the house two birds, purificabit eam reddita sanitate 14:49 et in purificationem

and cedar wood, and scarlet, and hyssop:

eius sumet duos passeres lignumque cedrinum et uermicu-

14:50 And he shall kill the one of the birds in an earthen lum atque hysopum 14:50 et immolato uno passere in uase vessel over running water:

14:51 And he shall take the cedar wood, and the hyssop, fictili super aquas uiuas 14:51 tollet lignum cedrinum et hy-

and the scarlet, and the living bird, and dip them in the blood of the slain bird, and in the running water, and sprin-

kle the house seven times: sopum et coccum et passerem uiuum et intinguet omnia in

sanguine passeris immolati atque in aquis uiuentibus et as-

14:52 And he shall cleanse the house with the blood of the perget domum septies 14:52 purificabitque eam tam in san-

bird, and with the running water, and with the living bird, and with the cedar wood, and with the hyssop, and with

the scarlet: guine passeris quam in aquis uiuentibus et in passere uiuo

14:53 But he shall let go the living bird out of the city into lignoque cedrino et hysopo atque uermiculo 14:53 cumque the open fields, and make an atonement for the house: and it shall be clean.

Liber Leuiticus

dimiserit passerem auolare in agrum libere orabit pro domo et iure mundabitur 14:54 ista est lex omnis leprae et percus- 14:54 This is the law for all manner of plague of leprosy,

and scall,

surae 14:55 leprae uestium et domorum 14:56 cicatricis et 14:55 And for the leprosy of a garment, and of a house,

14:56 And for a rising, and for a scab, and for a bright spot:

erumpentium papularum lucentis maculae et in uarias species coloribus inmutatis 14:57 ut possit sciri quo tempore 14:57 To teach when it is unclean, and when it is clean:

this is the law of leprosy.

mundum quid uel inmundum sit

15:1 locutusque est Dominus ad Mosen et Aaron dicens 15:1 And the LORD spake unto Moses and to Aaron, say-

ing,

15:2 loquimini filiis Israhel et dicite eis uir qui patitur fluxum 15:2 Speak unto the children of Israel, and say unto them,

When any man hath a running issue out of his flesh, because of his issue he is unclean.

seminis inmundus erit 15:3 et tunc iudicabitur huic uitio su-

15:3 And this shall be his uncleanness in his issue: whether his flesh run with his issue, or his flesh be stopped from his issue, it is his uncleanness.

biacere cum per momenta singula adheserit carni illius atque concreuerit foedus humor 15:4 omne stratum in quo dor- 15:4 Every bed, whereon he lieth that hath the issue, is

unclean: and every thing, whereon he sitteth, shall be unclean.

mierit inmundum erit et ubicumque sederit 15:5 si quis ho-

15:5 And whosoever toucheth his bed shall wash his clothes, and bathe himself in water, and be unclean until the even.

minum tetigerit lectum eius lauabit uestimenta sua et ipse lotus aqua inmundus erit usque ad uesperum 15:6 si sederit 15:6 And he that sitteth on any thing whereon he sat that

hath the issue shall wash his clothes, and bathe himself in water, and be unclean until the even.

ubi ille sederat et ipse lauabit uestimenta sua et lotus aqua inmundus erit usque ad uesperum 15:7 qui tetigerit carnem 15:7 And he that toucheth the flesh of him that hath the

issue shall wash his clothes, and bathe himself in water, and be unclean until the even.

eius lauabit uestimenta sua et ipse lotus aqua inmundus erit usque ad uesperum 15:8 si saliuam huiuscemodi homo ie- 15:8 And if he that hath the issue spit upon him that is

clean; then he shall wash his clothes, and bathe himself in water, and be unclean until the even.

cerit super eum qui mundus est lauabit uestem suam et lotus

aqua inmundus erit usque ad uesperum 15:9 sagma super 15:9 And what saddle soever he rideth upon that hath the

issue shall be unclean.

quo sederit inmundum erit 15:10 et quicquid sub eo fue- 15:10 And whosoever toucheth any thing that was under

him shall be unclean until the even: and he that beareth any of those things shall wash his clothes, and bathe himself

rit qui fluxum seminis patitur pollutum erit usque ad ues- in water, and be unclean until the even. perum qui portauerit horum aliquid lauabit uestem suam et ipse lotus aqua inmundus erit usque ad uesperum 15:11 om- 15:11 And whomsoever he toucheth that hath the issue,

and hath not rinsed his hands in water, he shall wash his clothes, and bathe himself in water, and be unclean until

nis quem tetigerit qui talis est non lotis ante manibus lauabit the even. uestimenta sua et lotus aqua inmundus erit usque ad uesperum 15:12 uas fictile quod tetigerit confringetur uas autem 15:12 And the vessel of earth, that he toucheth which hath

the issue, shall be broken: and every vessel of wood shall be rinsed in water.

ligneum lauabitur aqua 15:13 si sanatus fuerit qui huiusce-

15:13 And when he that hath an issue is cleansed of his issue; then he shall number to himself seven days for his cleansing, and wash his clothes, and bathe his flesh in running water, and shall be clean.

modi sustinet passionem numerabit septem dies post emundationem sui et lotis uestibus ac toto corpore in aquis uiuentibus erit mundus 15:14 die autem octauo sumet duos 15:14 And on the eighth day he shall take to him two

turtledoves, or two young pigeons, and come before the LORD unto the door of the tabernacle of the congregation,

turtures aut duos pullos columbae et ueniet in conspectu Do- and give them unto the priest: mini ad ostium tabernaculi testimonii dabitque eos sacerdoti

Biblia Sacra Vulgata

15:15 And the priest shall offer them, the one for a sin 15:15 qui faciet unum pro peccato et alterum in holocaus-

offering, and the other for a burnt offering; and the priest shall make an atonement for him before the LORD for his

issue. tum rogabitque pro eo coram Domino ut emundetur a fluxu 15:16 And if any man’s seed of copulation go out from seminis sui 15:16 uir de quo egreditur semen coitus lauabit

him, then he shall wash all his flesh in water, and be unclean until the even.

aqua omne corpus suum et inmundus erit usque ad uesperum

15:17 And every garment, and every skin, whereon is the 15:17 uestem et pellem quam habuerit lauabit aqua et in-

seed of copulation, shall be washed with water, and be unclean until the even.

munda erit usque ad uesperum 15:18 mulier cum qua coie-

15:18 The woman also with whom man shall lie with seed of copulation, they shall both bathe themselves in water, and be unclean until the even.

rit lauabitur aqua et inmunda erit usque ad uesperum 15:19

15:19 And if a woman have an issue, and her issue in her flesh be blood, she shall be put apart seven days: and

whosoever toucheth her shall be unclean until the even. mulier quae redeente mense patitur fluxum sanguinis septem 15:20 And every thing that she lieth upon in her separation diebus separabitur 15:20 omnis qui tetigerit eam inmundus

shall be unclean: every thing also that she sitteth upon shall be unclean.

erit usque ad uesperum 15:21 et in quo dormierit uel sederit

15:21 And whosoever toucheth her bed shall wash his clothes, and bathe himself in water, and be unclean until the even.

diebus separationis suae polluetur 15:22 qui tetigerit lectum

15:22 And whosoever toucheth any thing that she sat upon shall wash his clothes, and bathe himself in water, and be

unclean until the even. eius lauabit uestimenta sua et ipse lotus aqua inmundus erit 15:23 And if it be on her bed, or on any thing whereon she usque ad uesperum 15:23 omne uas super quo illa sederit

sitteth, when he toucheth it, he shall be unclean until the even.

quisquis adtigerit lauabit uestimenta sua et lotus aqua pollu-

15:24 And if any man lie with her at all, and her flowers tus erit usque ad uesperum 15:24 si coierit cum ea uir tem-

be upon him, he shall be unclean seven days; and all the bed whereon he lieth shall be unclean.

pore sanguinis menstrualis inmundus erit septem diebus et

15:25 And if a woman have an issue of her blood many omne stratum in quo dormierit polluetur 15:25 mulier quae

days out of the time of her separation, or if it run beyond the time of her separation; all the days of the issue of her uncleanness shall be as the days of her separation: she shall be unclean.

patitur multis diebus fluxum sanguinis non in tempore menstruali uel quae post menstruum sanguinem fluere non cessat quamdiu huic subiacet passioni inmunda erit quasi sit in

15:26 Every bed whereon she lieth all the days of her issue tempore menstruo 15:26 omne stratum in quo dormierit et shall be unto her as the bed of her separation: and whatsoever she sitteth upon shall be unclean, as the uncleanness

of her separation. uas in quo sederit pollutum erit 15:27 quicumque tetigerit

15:27 And whosoever toucheth those things shall be unclean, and shall wash his clothes, and bathe himself in water, and be unclean until the even.

eam lauabit uestimenta sua et ipse lotus aqua inmundus erit

15:28 But if she be cleansed of her issue, then she shall usque ad uesperum 15:28 si steterit sanguis et fluere ces-

number to herself seven days, and after that she shall be clean.

sarit numerabit septem dies purificationis suae 15:29 et oc-

15:29 And on the eighth day she shall take unto her two turtles, or two young pigeons, and bring them unto the priest, to the door of the tabernacle of the congregation.

tauo die offeret pro se sacerdoti duos turtures uel duos pullos

15:30 And the priest shall offer the one for a sin offering, columbae ad ostium tabernaculi testimonii 15:30 qui unum

and the other for a burnt offering; and the priest shall make an atonement for her before the LORD for the issue of her

uncleanness. faciet pro peccato et alterum in holocaustum rogabitque pro

15:31 Thus shall ye separate the children of Israel from ea coram Domino et pro fluxu inmunditiae eius 15:31 dotheir uncleanness; that they die not in their uncleanness, when they defile my tabernacle that is among them.

cebitis ergo filios Israhel ut caueant inmunditiam et non moriantur in sordibus suis cum polluerint tabernaculum meum

15:32 This is the law of him that hath an issue, and of him quod est inter eos 15:32 ista est lex eius qui patitur fluxum

whose seed goeth from him, and is defiled therewith;

15:33 And of her that is sick of her flowers, and of him seminis et qui polluitur coitu 15:33 et quae menstruis tem-

that hath an issue, of the man, and of the woman, and of him that lieth with her that is unclean.

Liber Leuiticus

poribus separatur uel quae iugi fluit sanguine et hominis qui dormierit cum ea

16:1 locutusque est Dominus ad Mosen post mortem duum 16:1 And the LORD spake unto Moses after the death

of the two sons of Aaron, when they offered before the LORD, and died;

filiorum Aaron quando offerentes ignem alienum interfecti sunt 16:2 et praecepit ei dicens loquere ad Aaron fratrem 16:2 And the LORD said unto Moses, Speak unto Aaron

thy brother, that he come not at all times into the holy place within the vail before the mercy seat, which is upon the ark; that he die not: for I will appear in the cloud upon the mercy seat.

tuum ne omni tempore ingrediatur sanctuarium quod est intra uelum coram propitiatorio quo tegitur arca ut non moriatur quia in nube apparebo super oraculum 16:3 nisi haec 16:3 Thus shall Aaron come into the holy place: with a

young bullock for a sin offering, and a ram for a burnt offering.

ante fecerit uitulum offeret pro peccato et arietem in holocaustum 16:4 tunica linea uestietur feminalibus lineis uere- 16:4 He shall put on the holy linen coat, and he shall have

the linen breeches upon his flesh, and shall be girded with a linen girdle, and with the linen mitre shall he be attired: these are holy garments; therefore shall he wash his flesh in water, and so put them on.

cunda celabit accingetur zona linea cidarim lineam inponet capiti haec enim uestimenta sunt sancta quibus cunctis cum lotus fuerit induetur 16:5 suscipietque ab uniuersa multitu- 16:5 And he shall take of the congregation of the children

of Israel two kids of the goats for a sin offering, and one ram for a burnt offering.

dine filiorum Israhel duos hircos pro peccato et unum arietem in holocaustum 16:6 cumque obtulerit uitulum et ora- 16:6 And Aaron shall offer his bullock of the sin offering,

which is for himself, and make an atonement for himself, and for his house.

uerit pro se et pro domo sua 16:7 duos hircos stare faciet

16:7 And he shall take the two goats, and present them before the LORD at the door of the tabernacle of the congregation.

coram Domino in ostio tabernaculi testimonii 16:8 mittens

16:8 And Aaron shall cast lots upon the two goats; one lot for the LORD, and the other lot for the scapegoat.

super utrumque sortem unam Domino et alteram capro emissario 16:9 cuius sors exierit Domino offeret illum pro pec- 16:9 And Aaron shall bring the goat upon which the

LORD’s lot fell, and offer him for a sin offering.

cato 16:10 cuius autem in caprum emissarium statuet eum 16:10 But the goat, on which the lot fell to be the scape-

goat, shall be presented alive before the LORD, to make an atonement with him, and to let him go for a scapegoat

uiuum coram Domino ut fundat preces super eo et emittat into the wilderness. illum in solitudinem 16:11 his rite celebratis offeret uitu- 16:11 And Aaron shall bring the bullock of the sin offer-

ing, which is for himself, and shall make an atonement for himself, and for his house, and shall kill the bullock of the

lum et rogans pro se et pro domo sua immolabit eum 16:12 sin offering which is for himself:

16:12 And he shall take a censer full of burning coals of fire from off the altar before the LORD, and his hands full of sweet incense beaten small, and bring it within the vail:

adsumptoque turibulo quod de prunis altaris impleuerit et hauriens manu conpositum thymiama in incensum ultra uelum intrabit in sancta 16:13 ut positis super ignem aromati- 16:13 And he shall put the incense upon the fire before the

LORD, that the cloud of the incense may cover the mercy seat that is upon the testimony, that he die not:

bus nebula eorum et uapor operiat oraculum quod est super testimonium et non moriatur 16:14 tollet quoque de san- 16:14 And he shall take of the blood of the bullock, and

sprinkle it with his finger upon the mercy seat eastward; and before the mercy seat shall he sprinkle of the blood

guine uituli et asperget digito septies contra propitiatorium with his finger seven times. ad orientem 16:15 cumque mactauerit hircum pro peccato 16:15 Then shall he kill the goat of the sin offering, that

is for the people, and bring his blood within the vail, and do with that blood as he did with the blood of the bullock, and sprinkle it upon the mercy seat, and before the mercy seat:

populi inferet sanguinem eius intra uelum sicut praeceptum est de sanguine uituli ut aspergat e regione oraculi 16:16 et 16:16 And he shall make an atonement for the holy place,

because of the uncleanness of the children of Israel, and because of their transgressions in all their sins: and so shall he do for the tabernacle of the congregation, that remaineth among them in the midst of their uncleanness.

expiet sanctuarium ab inmunditiis filiorum Israhel et a prae-

Biblia Sacra Vulgata

uaricationibus eorum cunctisque peccatis iuxta hunc ritum faciet tabernaculo testimonii quod fixum est inter eos in me-

16:17 And there shall be no man in the tabernacle of the dio sordium habitationis eorum 16:17 nullus hominum sit

congregation when he goeth in to make an atonement in the holy place, until he come out, and have made an atonement for himself, and for his household, and for all the congregation of Israel.

in tabernaculo quando pontifex ingreditur sanctuarium ut roget pro se et pro domo sua et pro uniuerso coetu Israhel do-

16:18 And he shall go out unto the altar that is before the nec egrediatur 16:18 cum autem exierit ad altare quod co-

LORD, and make an atonement for it; and shall take of the blood of the bullock, and of the blood of the goat, and put

it upon the horns of the altar round about. ram Domino est oret pro se et sumptum sanguinem uituli at-16:19 And he shall sprinkle of the blood upon it with his que hirci fundat super cornua eius per gyrum 16:19 asper-

finger seven times, and cleanse it, and hallow it from the uncleanness of the children of Israel.

gensque digito septies expiet et sanctificet illud ab inmun-

16:20 And when he hath made an end of reconciling the ditiis filiorum Israhel 16:20 postquam emundarit sanctua-

holy place, and the tabernacle of the congregation, and the altar, he shall bring the live goat:

rium et tabernaculum et altare tunc offerat hircum uiuen-

16:21 And Aaron shall lay both his hands upon the head of tem 16:21 et posita utraque manu super caput eius confithe live goat, and confess over him all the iniquities of the children of Israel, and all their transgressions in all their sins, putting them upon the head of the goat, and shall send him away by the hand of a fit man into the wilderness:

teatur omnes iniquitates filiorum Israhel et uniuersa delicta atque peccata eorum quae inprecans capiti eius emittet illum

16:22 And the goat shall bear upon him all their iniquities per hominem paratum in desertum 16:22 cumque portaueunto a land not inhabited: and he shall let go the goat in the wilderness.

rit hircus omnes iniquitates eorum in terram solitariam et di-

16:23 And Aaron shall come into the tabernacle of the missus fuerit in deserto 16:23 reuertetur Aaron in taberna-

congregation, and shall put off the linen garments, which he put on when he went into the holy place, and shall leave

them there: culum testimonii et depositis uestibus quibus prius indutus 16:24 And he shall wash his flesh with water in the holy erat cum intraret sanctuarium relictisque ibi 16:24 lauabit

place, and put on his garments, and come forth, and offer his burnt offering, and the burnt offering of the people, and

make an atonement for himself, and for the people. carnem suam in loco sancto indueturque uestimentis suis et

postquam egressus obtulerit holocaustum suum ac plebis ro-

16:25 And the fat of the sin offering shall he burn upon the gabit tam pro se quam pro populo 16:25 et adipem qui oblaaltar.

16:26 And he that let go the goat for the scapegoat shall tus est pro peccatis adolebit super altare 16:26 ille uero qui

wash his clothes, and bathe his flesh in water, and afterward come into the camp.

dimiserit caprum emissarium lauabit uestimenta sua et cor-

16:27 And the bullock for the sin offering, and the goat pus aqua et sic ingredietur in castra 16:27 uitulum autem

for the sin offering, whose blood was brought in to make atonement in the holy place, shall one carry forth without the camp; and they shall burn in the fire their skins, and their flesh, and their dung.

et hircum qui pro peccato fuerant immolati et quorum sanguis inlatus est ut in sanctuario expiatio conpleretur asportabunt foras castra et conburent igni tam pelles quam carnes

16:28 And he that burneth them shall wash his clothes, and eorum et fimum 16:28 et quicumque conbuserit ea lauabit bathe his flesh in water, and afterward he shall come into the camp.

uestimenta sua et carnem aqua et sic ingredietur in castra

16:29 And this shall be a statute for ever unto you: that in 16:29 eritque hoc uobis legitimum sempiternum mense sep-

the seventh month, on the tenth day of the month, ye shall afflict your souls, and do no work at all, whether it be one of your own country, or a stranger that sojourneth among you:

timo decima die mensis adfligetis animas uestras nullumque facietis opus siue indigena siue aduena qui peregrinatur in-

Liber Leuiticus

ter uos 16:30 in hac die expiatio erit uestri atque munda- 16:30 For on that day shall the priest make an atonement

for you, to cleanse you, that ye may be clean from all your sins before the LORD.

tio ab omnibus peccatis uestris coram Domino mundabimini 16:31 sabbatum enim requietionis est et adfligetis animas ue- 16:31 It shall be a sabbath of rest unto you, and ye shall

afflict your souls, by a statute for ever.

stras religione perpetua 16:32 expiabit autem sacerdos qui 16:32 And the priest, whom he shall anoint, and whom

he shall consecrate to minister in the priest’s office in his father’s stead, shall make the atonement, and shall put on

unctus fuerit et cuius initiatae manus ut sacerdotio funga- the linen clothes, even the holy garments: tur pro patre suo indueturque stola linea et uestibus sanctis 16:33 et expiabit sanctuarium et tabernaculum testimonii at- 16:33 And he shall make an atonement for the holy sanc-

tuary, and he shall make an atonement for the tabernacle of the congregation, and for the altar, and he shall make an atonement for the priests, and for all the people of the congregation.

que altare sacerdotes quoque et uniuersum populum 16:34

16:34 And this shall be an everlasting statute unto you, to make an atonement for the children of Israel for all their sins once a year. And he did as the LORD commanded Moses.

eritque hoc uobis legitimum sempiternum ut oretis pro filiis Israhel et pro cunctis peccatis eorum semel in anno fecit igitur sicut praeceperat Dominus Mosi

17:1 et locutus est Dominus ad Mosen dicens 17:2 lo- 17:1 And the LORD spake unto Moses, saying,

17:2 Speak unto Aaron, and unto his sons, and unto all the children of Israel, and say unto them; This is the thing which the LORD hath commanded, saying,

quere Aaron et filiis eius et cunctis filiis Israhel et dices

ad eos iste est sermo quem mandauit Dominus dicens 17:3 17:3 What man soever there be of the house of Israel, that killeth an ox, or lamb, or goat, in the camp, or that killeth it out of the camp,

homo quilibet de domo Israhel si occiderit bouem aut ouem siue capram in castris uel extra castra 17:4 et non obtule- 17:4 And bringeth it not unto the door of the tabernacle

of the congregation, to offer an offering unto the LORD before the tabernacle of the LORD; blood shall be imputed unto that man; he hath shed blood; and that man shall be cut off from among his people:

rit ad ostium tabernaculi oblationem Domino sanguinis reus erit quasi sanguinem fuderit sic peribit de medio populi sui 17:5 ideo offerre debent sacerdoti filii Israhel hostias suas 17:5 To the end that the children of Israel may bring their

sacrifices, which they offer in the open field, even that they may bring them unto the LORD, unto the door of the tabernacle of the congregation, unto the priest, and offer them for peace offerings unto the LORD.

quas occidunt in agro ut sanctificentur Domino ante ostium tabernaculi testimonii et immolent eas hostias pacificas Do-

mino 17:6 fundetque sacerdos sanguinem super altare Do- 17:6 And the priest shall sprinkle the blood upon the altar of the LORD at the door of the tabernacle of the congregation, and burn the fat for a sweet savour unto the LORD.

mini ad ostium tabernaculi testimonii et adolebit adipem in

odorem suauitatis Domino 17:7 et nequaquam ultra immo- 17:7 And they shall no more offer their sacrifices unto devils, after whom they have gone a whoring. This shall be a statute for ever unto them throughout their generations.

labunt hostias suas daemonibus cum quibus fornicati sunt legitimum sempiternum erit illis et posteris eorum 17:8 et 17:8 And thou shalt say unto them, Whatsoever man there

be of the house of Israel, or of the strangers which sojourn among you, that offereth a burnt offering or sacrifice,

ad ipsos dices homo de domo Israhel et de aduenis qui peregrinantur apud uos qui obtulerit holocaustum siue uictimam 17:9 et ad ostium tabernaculi testimonii non adduxerit eam ut 17:9 And bringeth it not unto the door of the tabernacle of

the congregation, to offer it unto the LORD; even that man shall be cut off from among his people.

offeratur Domino interibit de populo suo 17:10 homo qui-

17:10 And whatsoever man there be of the house of Israel, or of the strangers that sojourn among you, that eateth any manner of blood; I will even set my face against that soul that eateth blood, and will cut him off from among his people.

libet de domo Israhel et de aduenis qui peregrinantur inter eos si comederit sanguinem obfirmabo faciem meam contra animam illius et disperdam eam de populo suo 17:11 quia 17:11 For the life of the flesh is in the blood: and I have

given it to you upon the altar to make an atonement for your souls: for it is the blood that maketh an atonement for the soul.

Biblia Sacra Vulgata

anima carnis in sanguine est et ego dedi illum uobis ut super altare in eo expietis pro animabus uestris et sanguis pro ani-

17:12 Therefore I said unto the children of Israel, No soul mae piaculo sit 17:12 idcirco dixi filiis Israhel omnis anima

of you shall eat blood, neither shall any stranger that sojourneth among you eat blood.

ex uobis non comedet sanguinem nec ex aduenis qui pereg-

17:13 And whatsoever man there be of the children of Is- rinantur inter uos 17:13 homo quicumque de filiis Israhel

rael, or of the strangers that sojourn among you, which hunteth and catcheth any beast or fowl that may be eaten; he shall even pour out the blood thereof, and cover it with dust.

et de aduenis qui peregrinantur apud uos si uenatione atque aucupio ceperit feram uel auem quibus uesci licitum est fun-

17:14 For it is the life of all flesh; the blood of it is for the dat sanguinem eius et operiat illum terra 17:14 anima enim life thereof: therefore I said unto the children of Israel, Ye shall eat the blood of no manner of flesh: for the life of all flesh is the blood thereof: whosoever eateth it shall be cut off.

omnis carnis in sanguine est unde dixi filiis Israhel sanguinem uniuersae carnis non comedetis quia anima carnis in

17:15 And every soul that eateth that which died of itself, sanguine est et quicumque comederit illum interibit 17:15

or that which was torn with beasts, whether it be one of your own country, or a stranger, he shall both wash his clothes, and bathe himself in water, and be unclean until the even: then shall he be clean.

anima quae comederit morticinum uel captum a bestia tam de indigenis quam de aduenis lauabit uestes suas et semet ipsum aqua et contaminatus erit usque ad uesperum et hoc

17:16 But if he wash them not, nor bathe his flesh; then he ordine mundus fiet 17:16 quod si non lauerit uestimenta sua

shall bear his iniquity.

nec corpus portabit iniquitatem suam

18:1 And the LORD spake unto Moses, saying, 18:1 locutusque est Dominus ad Mosen dicens 18:2 lo-

18:2 Speak unto the children of Israel, and say unto them, I am the LORD your God.

quere filiis Israhel et dices ad eos ego Dominus Deus uester

18:3 After the doings of the land of Egypt, wherein ye 18:3 iuxta consuetudinem terrae Aegypti in qua habitastis

dwelt, shall ye not do: and after the doings of the land of Canaan, whither I bring you, shall ye not do: neither shall

ye walk in their ordinances. non facietis et iuxta morem regionis Chanaan ad quam ego

introducturus sum uos non agetis nec in legitimis eorum am-

18:4 Ye shall do my judgments, and keep mine ordinances, bulabitis 18:4 facietis iudicia mea et praecepta seruabitis et

to walk therein: I am the LORD your God.

18:5 Ye shall therefore keep my statutes, and my judg- ambulabitis in eis ego Dominus Deus uester 18:5 custodite

ments: which if a man do, he shall live in them: I am the LORD.

leges meas atque iudicia quae faciens homo uiuet in eis ego

18:6 None of you shall approach to any that is near of kin Dominus 18:6 omnis homo ad proximam sanguinis sui non to him, to uncover their nakedness: I am the LORD.

18:7 The nakedness of thy father, or the nakedness of thy accedet ut reuelet turpitudinem eius ego Dominus 18:7 tur-

mother, shalt thou not uncover: she is thy mother; thou shalt not uncover her nakedness.

pitudinem patris et turpitudinem matris tuae non discoperies

18:8 The nakedness of thy father’s wife shalt thou not un- mater tua est non reuelabis turpitudinem eius 18:8 turpitu-

cover: it is thy father’s nakedness.

dinem uxoris patris tui non discoperies turpitudo enim patris

18:9 The nakedness of thy sister, the daughter of thy fa- tui est 18:9 turpitudinem sororis tuae ex patre siue ex matre

ther, or daughter of thy mother, whether she be born at home, or born abroad, even their nakedness thou shalt not

uncover. quae domi uel foris genita est non reuelabis 18:10 turpi-

18:10 The nakedness of thy son’s daughter, or of thy daughter’s daughter, even their nakedness thou shalt not uncover: for theirs is thine own nakedness.

tudinem filiae filii tui uel neptis ex filia non reuelabis quia

18:11 The nakedness of thy father’s wife’s daughter, be- turpitudo tua est 18:11 turpitudinem filiae uxoris patris tui

gotten of thy father, she is thy sister, thou shalt not uncover her nakedness.

Liber Leuiticus

quam peperit patri tuo et est soror tua non reuelabis 18:12 18:12 Thou shalt not uncover the nakedness of thy father’s

sister: she is thy father’s near kinswoman.

turpitudinem sororis patris tui non discoperies quia caro est patris tui 18:13 turpitudinem sororis matris tuae non reue- 18:13 Thou shalt not uncover the nakedness of thy

mother’s sister: for she is thy mother’s near kinswoman.

labis eo quod caro sit matris tuae 18:14 turpitudinem patrui 18:14 Thou shalt not uncover the nakedness of thy father’s

brother, thou shalt not approach to his wife: she is thine aunt.

tui non reuelabis nec accedes ad uxorem eius quae tibi adfinitate coniungitur 18:15 turpitudinem nurus tuae non reue- 18:15 Thou shalt not uncover the nakedness of thy daugh-

ter in law: she is thy son’s wife; thou shalt not uncover her nakedness.

labis quia uxor filii tui est nec discoperies ignominiam eius 18:16 turpitudinem uxoris fratris tui non reuelabis quia tur- 18:16 Thou shalt not uncover the nakedness of thy

brother’s wife: it is thy brother’s nakedness.

pitudo fratris tui est 18:17 turpitudinem uxoris tuae et filiae 18:17 Thou shalt not uncover the nakedness of a woman

and her daughter, neither shalt thou take her son’s daughter, or her daughter’s daughter, to uncover her nakedness;

eius non reuelabis filiam filii eius et filiam filiae illius non su- for they are her near kinswomen: it is wickedness. mes ut reueles ignominiam eius quia caro illius sunt et talis coitus incestus est 18:18 sororem uxoris tuae in pelicatum 18:18 Neither shalt thou take a wife to her sister, to vex

her, to uncover her nakedness, beside the other in her life time.

illius non accipies nec reuelabis turpitudinem eius adhuc illa uiuente 18:19 ad mulierem quae patitur menstrua non ac- 18:19 Also thou shalt not approach unto a woman to un-

cover her nakedness, as long as she is put apart for her uncleanness.

cedes nec reuelabis foeditatem eius 18:20 cum uxore pro-

18:20 Moreover thou shalt not lie carnally with thy neighbour’s wife, to defile thyself with her.

ximi tui non coibis nec seminis commixtione maculaberis 18:21 de semine tuo non dabis ut consecretur idolo Moloch 18:21 And thou shalt not let any of thy seed pass through

the fire to Molech, neither shalt thou profane the name of thy God: I am the LORD.

nec pollues nomen Dei tui ego Dominus 18:22 cum mas-

18:22 Thou shalt not lie with mankind, as with womankind: it is abomination.

culo non commisceberis coitu femineo quia abominatio est

18:23 cum omni pecore non coibis nec maculaberis cum eo 18:23 Neither shalt thou lie with any beast to defile thyself therewith: neither shall any woman stand before a beast to lie down thereto: it is confusion.

mulier non subcumbet iumento nec miscebitur ei quia scelus est 18:24 ne polluamini in omnibus his quibus contaminatae 18:24 Defile not ye yourselves in any of these things: for

in all these the nations are defiled which I cast out before you:

sunt uniuersae gentes quas ego eiciam ante conspectum uestrum 18:25 et quibus polluta est terra cuius ego scelera ui- 18:25 And the land is defiled: therefore I do visit the in-

iquity thereof upon it, and the land itself vomiteth out her inhabitants.

sitabo ut euomat habitatores suos 18:26 custodite legitima

18:26 Ye shall therefore keep my statutes and my judgments, and shall not commit any of these abominations; neither any of your own nation, nor any stranger that sojourneth among you:

mea atque iudicia et non faciat ex omnibus abominationibus istis tam indigena quam colonus qui peregrinatur apud uos 18:27 omnes enim execrationes istas fecerunt accolae terrae 18:27 (For all these abominations have the men of the land

done, which were before you, and the land is defiled;)

qui fuerunt ante uos et polluerunt eam 18:28 cauete ergo 18:28 That the land spue not you out also, when ye defile

it, as it spued out the nations that were before you.

ne et uos similiter euomat cum paria feceritis sicut euomuit gentem quae fuit ante uos 18:29 omnis anima quae fecerit 18:29 For whosoever shall commit any of these abomi-

nations, even the souls that commit them shall be cut off from among their people.

de abominationibus his quippiam peribit de medio populi sui 18:30 custodite mandata mea nolite facere quae fecerunt hii 18:30 Therefore shall ye keep mine ordinance, that ye

commit not any one of these abominable customs, which were committed before you, and that ye defile not yourselves therein: I am the LORD your God.

Biblia Sacra Vulgata

qui fuerunt ante uos et ne polluamini in eis ego Dominus Deus uester

19:1 And the LORD spake unto Moses, saying, 19:1 locutus est Dominus ad Mosen dicens 19:2 loquere

19:2 Speak unto all the congregation of the children of Israel, and say unto them, Ye shall be holy: for I the LORD your God am holy.

ad omnem coetum filiorum Israhel et dices ad eos sancti es-

19:3 Ye shall fear every man his mother, and his father, tote quia ego sanctus sum Dominus Deus uester 19:3 unus-

and keep my sabbaths: I am the LORD your God.

quisque matrem et patrem suum timeat sabbata mea custo-

19:4 Turn ye not unto idols, nor make to yourselves molten dite ego Dominus Deus uester 19:4 nolite conuerti ad idola

gods: I am the LORD your God.

nec deos conflatiles faciatis uobis ego Dominus Deus ues-

19:5 And if ye offer a sacrifice of peace offerings unto the ter 19:5 si immolaueritis hostiam pacificorum Domino ut

LORD, ye shall offer it at your own will.

19:6 It shall be eaten the same day ye offer it, and on the sit placabilis 19:6 eo die quo fuerit immolata comedetis

morrow: and if ought remain until the third day, it shall be burnt in the fire.

eam et die altero quicquid autem residuum fuerit in diem ter-

19:7 And if it be eaten at all on the third day, it is abom- tium igne conburetis 19:7 si quis post biduum comederit ex

inable; it shall not be accepted.

19:8 Therefore every one that eateth it shall bear his iniq- ea profanus erit et impietatis reus 19:8 portabit iniquitatem

uity, because he hath profaned the hallowed thing of the LORD: and that soul shall be cut off from among his peo-

ple. suam quia sanctum Domini polluit et peribit anima illa de 19:9 And when ye reap the harvest of your land, thou shalt populo suo 19:9 cum messueris segetes terrae tuae non ton-

not wholly reap the corners of thy field, neither shalt thou gather the gleanings of thy harvest.

debis usque ad solum superficiem terrae nec remanentes spi-

19:10 And thou shalt not glean thy vineyard, neither shalt cas colliges 19:10 neque in uinea tua racemos et grana deci-

thou gather every grape of thy vineyard; thou shalt leave them for the poor and stranger: I am the LORD your God.

dentia congregabis sed pauperibus et peregrinis carpenda di-

19:11 Ye shall not steal, neither deal falsely, neither lie one mittes ego Dominus Deus uester 19:11 non facietis furtum to another.

non mentiemini nec decipiet unusquisque proximum suum

19:12 And ye shall not swear by my name falsely, neither 19:12 non peierabis in nomine meo nec pollues nomen Dei tui

shalt thou profane the name of thy God: I am the LORD. 19:13 Thou shalt not defraud thy neighbour, neither rob ego Dominus 19:13 non facies calumniam proximo tuo nec him: the wages of him that is hired shall not abide with thee all night until the morning.

ui opprimes eum non morabitur opus mercennarii apud te us-

19:14 Thou shalt not curse the deaf, nor put a stumbling- que mane 19:14 non maledices surdo nec coram caeco poblock before the blind, but shalt fear thy God: I am the LORD.

nes offendiculum sed timebis Deum tuum quia ego sum Do-

19:15 Ye shall do no unrighteousness in judgment: thou minus 19:15 non facies quod iniquum est nec iniuste iudi-

shalt not respect the person of the poor, nor honor the person of the mighty: but in righteousness shalt thou judge

thy neighbour. cabis nec consideres personam pauperis nec honores uultum 19:16 Thou shalt not go up and down as a talebearer potentis iuste iudica proximo tuo 19:16 non eris criminator

among thy people: neither shalt thou stand against the blood of thy neighbour; I am the LORD.

et susurro in populis non stabis contra sanguinem proximi

19:17 Thou shalt not hate thy brother in thine heart: thou tui ego Dominus 19:17 ne oderis fratrem tuum in corde tuo shalt in any wise rebuke thy neighbour, and not suffer sin upon him.

sed publice argue eum ne habeas super illo peccatum 19:18

19:18 Thou shalt not avenge, nor bear any grudge against the children of thy people, but thou shalt love thy neighbour as thyself: I am the LORD.

non quaeres ultionem nec memor eris iniuriae ciuium tuorum diliges amicum tuum sicut temet ipsum ego Dominus

Liber Leuiticus

19:19 leges meas custodite iumenta tua non facies coire cum 19:19 Ye shall keep my statutes. Thou shalt not let thy

cattle gender with a diverse kind: thou shalt not sow thy field with mingled seed: neither shall a garment mingled

alterius generis animantibus agrum non seres diuerso semine of linen and woollen come upon thee. ueste quae ex duobus texta est non indueris 19:20 homo si 19:20 And whosoever lieth carnally with a woman, that is

a bondmaid, betrothed to an husband, and not at all redeemed, nor freedom given her; she shall be scourged;

dormierit cum muliere coitu seminis quae sit ancilla etiam they shall not be put to death, because she was not free. nubilis et tamen pretio non redeepta nec libertate donata uapulabunt ambo et non morientur quia non fuit libera 19:21 19:21 And he shall bring his trespass offering unto the

LORD, unto the door of the tabernacle of the congregation, even a ram for a trespass offering.

pro delicto autem suo offeret Domino ad ostium tabernaculi testimonii arietem 19:22 orabitque pro eo sacerdos et 19:22 And the priest shall make an atonement for him with

the ram of the trespass offering before the LORD for his sin which he hath done: and the sin which he hath done

pro delicto eius coram Domino et repropitiabitur ei dimit- shall be forgiven him. teturque peccatum 19:23 quando ingressi fueritis terram et 19:23 And when ye shall come into the land, and shall

have planted all manner of trees for food, then ye shall count the fruit thereof as uncircumcised: three years shall

plantaueritis in ea ligna pomifera auferetis praeputia eorum it be as uncircumcised unto you: it shall not be eaten of. poma quae germinant inmunda erunt uobis nec edetis ex eis 19:24 quarto anno omnis fructus eorum sanctificabitur lau- 19:24 But in the fourth year all the fruit thereof shall be

holy to praise the LORD withal.

dabilis Domino 19:25 quinto autem anno comedetis fructus 19:25 And in the fifth year shall ye eat of the fruit thereof,

that it may yield unto you the increase thereof: I am the LORD your God.

congregantes poma quae proferunt ego Dominus Deus uester 19:26 non comedetis cum sanguine non augurabimini 19:26 Ye shall not eat any thing with the blood: neither

shall ye use enchantment, nor observe times.

nec obseruabitis somnia 19:27 neque in rotundum adton- 19:27 Ye shall not round the corners of your heads, neither

shalt thou mar the corners of thy beard.

debitis comam nec radatis barbam 19:28 et super mortuo 19:28 Ye shall not make any cuttings in your flesh for the

dead, nor print any marks upon you: I am the LORD.

non incidetis carnem uestram neque figuras aliquas et stigmata facietis uobis ego Dominus 19:29 ne prostituas filiam 19:29 Do not prostitute thy daughter, to cause her to be

a whore; lest the land fall to whoredom, and the land become full of wickedness.

tuam et contaminetur terra et impleatur piaculo 19:30 sab-

19:30 Ye shall keep my sabbaths, and reverence my sanctuary: I am the LORD.

bata mea custodite et sanctuarium meum metuite ego Do-

minus 19:31 ne declinetis ad magos nec ab ariolis aliquid 19:31 Regard not them that have familiar spirits, neither seek after wizards, to be defiled by them: I am the LORD your God.

sciscitemini ut polluamini per eos ego Dominus Deus ues-

ter 19:32 coram cano capite consurge et honora personam 19:32 Thou shalt rise up before the hoary head, and honour

the face of the old man, and fear thy God: I am the LORD.

senis et time Deum tuum ego sum Dominus 19:33 si habi- 19:33 And if a stranger sojourn with thee in your land, ye

shall not vex him.

tauerit aduena in terra uestra et moratus fuerit inter uos ne exprobretis ei 19:34 sed sit inter uos quasi indigena et dili- 19:34 But the stranger that dwelleth with you shall be unto

you as one born among you, and thou shalt love him as thyself; for ye were strangers in the land of Egypt: I am

getis eum quasi uosmet ipsos fuistis enim et uos aduenae in the LORD your God. terra Aegypti ego Dominus Deus uester 19:35 nolite facere 19:35 Ye shall do no unrighteousness in judgment, in

meteyard, in weight, or in measure.

iniquum aliquid in iudicio in regula in pondere in mensura 19:36 statera iusta et aequa sint pondera iustus modius ae- 19:36 Just balances, just weights, a just ephah, and a just

hin, shall ye have: I am the LORD your God, which brought you out of the land of Egypt.

quusque sextarius ego Dominus Deus uester qui eduxi uos

Biblia Sacra Vulgata

19:37 Therefore shall ye observe all my statutes, and all de terra Aegypti 19:37 custodite omnia praecepta mea et my judgments, and do them: I am the LORD.

uniuersa iudicia et facite ea ego Dominus

20:1 And the LORD spake unto Moses, saying, 20:1 locutusque est Dominus ad Mosen dicens 20:2 haec

20:2 Again, thou shalt say to the children of Israel, Whosoever he be of the children of Israel, or of the strangers that sojourn in Israel, that giveth any of his seed unto Molech; he shall surely be put to death: the people of the land shall stone him with stones.

loqueris filiis Israhel homo de filiis Israhel et de aduenis qui habitant in Israhel si quis dederit de semine suo idolo Mo-

20:3 And I will set my face against that man, and will loch morte moriatur populus terrae lapidabit eum 20:3 et

cut him off from among his people; because he hath given of his seed unto Molech, to defile my sanctuary, and to

profane my holy name. ego ponam faciem meam contra illum succidamque eum de medio populi sui eo quod dederit de semine suo Moloch et contaminauerit sanctuarium meum ac polluerit nomen sanc-

20:4 And if the people of the land do any ways hide tum meum 20:4 quod si neglegens populus terrae et quasi

their eyes from the man, when he giveth of his seed unto Molech, and kill him not:

paruipendens imperium meum dimiserit hominem qui dede-

20:5 Then I will set my face against that man, and against rit de semine suo Moloch nec uoluerit eum occidere 20:5 his family, and will cut him off, and all that go a whoring after him, to commit whoredom with Molech, from among

their people. ponam faciem meam super hominem illum et cognationem

eius succidamque et ipsum et omnes qui consenserunt ei ut

20:6 And the soul that turneth after such as have familiar fornicarentur cum Moloch de medio populi sui 20:6 anima spirits, and after wizards, to go a whoring after them, I will even set my face against that soul, and will cut him

off from among his people. quae declinauerit ad magos et ariolos et fornicata fuerit cum

eis ponam faciem meam contra eam et interficiam illam de

20:7 Sanctify yourselves therefore, and be ye holy: for I medio populi sui 20:7 sanctificamini et estote sancti quia

am the LORD your God.

20:8 And ye shall keep my statutes, and do them: I am the ego Dominus Deus uester 20:8 custodite praecepta mea et

LORD which sanctify you.

20:9 For every one that curseth his father or his mother facite ea ego Dominus qui sanctifico uos 20:9 qui maledixe-

shall be surely put to death: he hath cursed his father or his mother; his blood shall be upon him.

rit patri suo et matri morte moriatur patri matrique maledixit

20:10 And the man that committeth adultery with another sanguis eius sit super eum 20:10 si moechatus quis fuerit

man’s wife, even he that committeth adultery with his neighbour’s wife, the adulterer and the adulteress shall

surely be put to death. cum uxore alterius et adulterium perpetrarit cum coniuge

20:11 And the man that lieth with his father’s wife hath un- proximi sui morte moriantur et moechus et adultera 20:11 covered his father’s nakedness: both of them shall surely be put to death; their blood shall be upon them.

qui dormierit cum nouerca sua et reuelauerit ignominiam patris sui morte moriantur ambo sanguis eorum sit super eos

20:12 And if a man lie with his daughter in law, both of 20:12 si quis dormierit cum nuru sua uterque moriantur quia

them shall surely be put to death: they have wrought confusion; their blood shall be upon them.

scelus operati sunt sanguis eorum sit super eos 20:13 qui

20:13 If a man also lie with mankind, as he lieth with a woman, both of them have committed an abomination: they shall surely be put to death; their blood shall be upon them.

dormierit cum masculo coitu femineo uterque operati sunt

20:14 And if a man take a wife and her mother, it is nefas morte moriantur sit sanguis eorum super eos 20:14

wickedness: they shall be burnt with fire, both he and they; that there be no wickedness among you.

qui supra uxorem filiam duxerit matrem eius scelus operatus est uiuus ardebit cum eis nec permanebit tantum nefas in

20:15 And if a man lie with a beast, he shall surely be put medio uestri 20:15 qui cum iumento et pecore coierit morte

to death: and ye shall slay the beast.

Liber Leuiticus

moriatur pecus quoque occidite 20:16 mulier quae subcu- 20:16 And if a woman approach unto any beast, and lie

down thereto, thou shalt kill the woman, and the beast: they shall surely be put to death; their blood shall be upon

buerit cuilibet iumento simul interficietur cum eo sanguis eo- them.

rum sit super eos 20:17 qui acceperit sororem suam filiam 20:17 And if a man shall take his sister, his father’s daughter, or his mother’s daughter, and see her nakedness, and she see his nakedness; it is a wicked thing; and they shall be cut off in the sight of their people: he hath uncovered his sister’s nakedness; he shall bear his iniquity.

patris sui uel filiam matris suae et uiderit turpitudinem eius illaque conspexerit fratris ignominiam nefariam rem operati sunt occidentur in conspectu populi sui eo quod turpitudinem suam mutuo reuelarint et portabunt iniquitatem suam 20:18 qui coierit cum muliere in fluxu menstruo et reuelaue- 20:18 And if a man shall lie with a woman having her sick-

ness, and shall uncover her nakedness; he hath discovered her fountain, and she hath uncovered the fountain of her blood: and both of them shall be cut off from among their people.

rit turpitudinem eius ipsaque aperuerit fontem sanguinis sui interficientur ambo de medio populi sui 20:19 turpitudinem 20:19 And thou shalt not uncover the nakedness of thy

mother’s sister, nor of thy father’s sister: for he uncovereth his near kin: they shall bear their iniquity.

materterae tuae et amitae tuae non discoperies qui hoc fecerit ignominiam carnis suae nudauit portabunt ambo iniquitatem suam 20:20 qui coierit cum uxore patrui uel auunculi sui 20:20 And if a man shall lie with his uncle’s wife, he hath

uncovered his uncle’s nakedness: they shall bear their sin; they shall die childless.

et reuelauerit ignominiam cognationis suae portabunt ambo peccatum suum absque liberis morientur 20:21 qui duxerit 20:21 And if a man shall take his brother’s wife, it is an

unclean thing: he hath uncovered his brother’s nakedness; they shall be childless.

uxorem fratris sui rem facit inlicitam turpitudinem fratris sui reuelauit absque filiis erunt 20:22 custodite leges meas at- 20:22 Ye shall therefore keep all my statutes, and all my

judgments, and do them: that the land, whither I bring you to dwell therein, spue you not out.

que iudicia et facite ea ne et uos euomat terra quam intraturi estis et habitaturi 20:23 nolite ambulare in legitimis natio- 20:23 And ye shall not walk in the manners of the nation,

which I cast out before you: for they committed all these things, and therefore I abhorred them.

num quas ego expulsurus sum ante uos omnia enim haec

fecerunt et abominatus sum eos 20:24 uobis autem loquor 20:24 But I have said unto you, Ye shall inherit their land, and I will give it unto you to possess it, a land that floweth with milk and honey: I am the LORD your God, which

possidete terram eorum quam dabo uobis in hereditatem ter- have separated you from other people. ram fluentem lacte et melle ego Dominus Deus uester qui separaui uos a ceteris populis 20:25 separate ergo et uos iu- 20:25 Ye shall therefore put difference between clean

beasts and unclean, and between unclean fowls and clean: and ye shall not make your souls abominable by beast, or by fowl, or by any manner of living thing that creepeth on the ground, which I have separated from you as unclean.

mentum mundum ab inmundo et auem mundam ab inmunda ne polluatis animas uestras in pecore et in auibus et cunctis quae mouentur in terra et quae uobis ostendi esse polluta 20:26 eritis sancti mihi quia sanctus ego sum Dominus et se- 20:26 And ye shall be holy unto me: for I the LORD

am holy, and have severed you from other people, that ye should be mine.

paraui uos a ceteris populis ut essetis mei 20:27 uir siue

20:27 A man also or woman that hath a familiar spirit, or that is a wizard, shall surely be put to death: they shall stone them with stones: their blood shall be upon them.

mulier in quibus pythonicus uel diuinationis fuerit spiritus morte moriantur lapidibus obruent eos sanguis eorum sit super illos

21:1 dixit quoque Dominus ad Mosen loquere ad sacer- 21:1 And the LORD said unto Moses, Speak unto the

priests the sons of Aaron, and say unto them, There shall none be defiled for the dead among his people:

Biblia Sacra Vulgata

dotes filios Aaron et dices eis ne contaminetur sacerdos in

21:2 But for his kin, that is near unto him, that is, for mortibus ciuium suorum 21:2 nisi tantum in consanguineis

his mother, and for his father, and for his son, and for his daughter, and for his brother.

ac propinquis id est super matre et patre et filio ac filia fratre

21:3 And for his sister a virgin, that is nigh unto him, quoque 21:3 et sorore uirgine quae non est nupta uiro 21:4

which hath had no husband; for her may he be defiled. 21:4 But he shall not defile himself, being a chief man among his people, to profane himself.

sed nec in principe populi sui contaminabitur 21:5 non ra-

21:5 They shall not make baldness upon their head, neither shall they shave off the corner of their beard, nor make any cuttings in their flesh.

dent caput nec barbam neque in carnibus suis facient incisu-

21:6 They shall be holy unto their God, and not profane the ras 21:6 sancti erunt Deo suo et non polluent nomen eius in-

name of their God: for the offerings of the LORD made by fire, and the bread of their God, they do offer: therefore

they shall be holy. censum enim Domini et panes Dei sui offerunt et ideo sancti 21:7 They shall not take a wife that is a whore, or pro- erunt 21:7 scortum et uile prostibulum non ducet uxorem

fane; neither shall they take a woman put away from her husband: for he is holy unto his God.

nec eam quae repudiata est a marito quia consecratus est

21:8 Thou shalt sanctify him therefore; for he offereth the Deo suo 21:8 et panes propositionis offert sit ergo sanctus

bread of thy God: he shall be holy unto thee: for I the LORD, which sanctify you, am holy.

quia et ego sanctus sum Dominus qui sanctifico uos 21:9

21:9 And the daughter of any priest, if she profane herself by playing the whore, she profaneth her father: she shall be burnt with fire.

sacerdotis filia si deprehensa fuerit in stupro et uiolauerit

21:10 And he that is the high priest among his brethren, nomen patris sui flammis exuretur 21:10 pontifex id est sa-

upon whose head the anointing oil was poured, and that is consecrated to put on the garments, shall not uncover his

head, nor rend his clothes; cerdos maximus inter fratres suos super cuius caput fusum est unctionis oleum et cuius manus in sacerdotio consecratae sunt uestitusque est sanctis uestibus caput suum non disco-

21:11 Neither shall he go in to any dead body, nor defile periet uestimenta non scindet 21:11 et ad omnem mortuum

himself for his father, or for his mother;

non ingredietur omnino super patre quoque suo et matre non

21:12 Neither shall he go out of the sanctuary, nor profane contaminabitur 21:12 nec egredietur de sanctis ne polluat

the sanctuary of his God; for the crown of the anointing oil of his God is upon him: I am the LORD.

sanctuarium Domini quia oleum sanctae unctionis Dei sui

21:13 And he shall take a wife in her virginity. super eum est ego Dominus 21:13 uirginem ducet uxorem 21:14 A widow, or a divorced woman, or profane, or an 21:14 uiduam et repudiatam et sordidam atque meretricem

harlot, these shall he not take: but he shall take a virgin of his own people to wife.

non accipiet sed puellam de populo suo 21:15 ne commi-

21:15 Neither shall he profane his seed among his people: for I the LORD do sanctify him.

sceat stirpem generis sui uulgo gentis suae quia ego Domi-

21:16 And the LORD spake unto Moses, saying, nus qui sanctifico eum 21:16 locutusque est Dominus ad

21:17 Speak unto Aaron, saying, Whosoever he be of thy Mosen dicens 21:17 loquere ad Aaron homo de semine tuo seed in their generations that hath any blemish, let him not approach to offer the bread of his God.

per familias qui habuerit maculam non offeret panes Deo

21:18 For whatsoever man he be that hath a blemish, he suo 21:18 nec accedet ad ministerium eius si caecus fue-

shall not approach: a blind man, or a lame, or he that hath a flat nose, or any thing superfluous,

rit si claudus si uel paruo uel grandi et torto naso 21:19 si

21:19 Or a man that is brokenfooted, or brokenhanded,

21:20 Or crookbackt, or a dwarf, or that hath a blemish fracto pede si manu 21:20 si gibbus si lippus si albuginem

in his eye, or be scurvy, or scabbed, or hath his stones broken;

habens in oculo si iugem scabiem si inpetiginem in corpore

21:21 No man that hath a blemish of the seed of Aaron the uel hirniosus 21:21 omnis qui habuerit maculam de semine

priest shall come nigh to offer the offerings of the LORD made by fire: he hath a blemish; he shall not come nigh to offer the bread of his God.

Liber Leuiticus

Aaron sacerdotis non accedet offerre hostias Domino nec panes Deo suo 21:22 uescetur tamen panibus qui offeruntur in 21:22 He shall eat the bread of his God, both of the most

holy, and of the holy.

sanctuario 21:23 ita dumtaxat ut intra uelum non ingredia- 21:23 Only he shall not go in unto the vail, nor come nigh

unto the altar, because he hath a blemish; that he profane not my sanctuaries: for I the LORD do sanctify them.

tur nec accedat ad altare quia maculam habet et contaminare non debet sanctuarium meum ego Dominus qui sanctifico eos 21:24 locutus est ergo Moses ad Aaron et filios eius et 21:24 And Moses told it unto Aaron, and to his sons, and

unto all the children of Israel.

ad omnem Israhel cuncta quae sibi fuerant imperata

22:1 locutus quoque est Dominus ad Mosen dicens 22:2 22:1 And the LORD spake unto Moses, saying,

22:2 Speak unto Aaron and to his sons, that they separate themselves from the holy things of the children of Israel, and that they profane not my holy name in those things which they hallow unto me: I am the LORD.

loquere ad Aaron et ad filios eius ut caueant ab his quae consecrata sunt filiorum Israhel et non contaminent nomen sanctificatorum mihi quae ipsi offerunt ego Dominus 22:3 dic 22:3 Say unto them, Whosoever he be of all your seed

among your generations, that goeth unto the holy things, which the children of Israel hallow unto the LORD, having his uncleanness upon him, that soul shall be cut off from my presence: I am the LORD.

ad eos et ad posteros eorum omnis homo qui accesserit de stirpe uestra ad ea quae consecrata sunt et quae obtulerunt filii Israhel Domino in quo est inmunditia peribit coram Do-

mino ego sum Dominus 22:4 homo de semine Aaron qui 22:4 What man soever of the seed of Aaron is a leper, or hath a running issue; he shall not eat of the holy things, until he be clean. And whoso toucheth any thing that is

fuerit leprosus aut patiens fluxum seminis non uescetur de unclean by the dead, or a man whose seed goeth from him; his quae sanctificata sunt mihi donec sanetur qui tetigerit inmundum super mortuo et ex quo egreditur semen quasi coitus 22:5 et qui tangit reptile et quodlibet inmundum cuius 22:5 Or whosoever toucheth any creeping thing, whereby

he may be made unclean, or a man of whom he may take uncleanness, whatsoever uncleanness he hath;

tactus est sordidus 22:6 inmundus erit usque ad uesperum

22:6 The soul which hath touched any such shall be unclean until even, and shall not eat of the holy things, unless he wash his flesh with water.

et non uescetur his quae sanctificata sunt sed cum lauerit carnem suam aqua 22:7 et occubuerit sol tunc mundatus ues- 22:7 And when the sun is down, he shall be clean, and

shall afterward eat of the holy things; because it is his food.

cetur de sanctificatis quia cibus illius est 22:8 morticinum

22:8 That which dieth of itself, or is torn with beasts, he shall not eat to defile himself therewith; I am the LORD.

et captum a bestia non comedent nec polluentur in eis ego

sum Dominus 22:9 custodient praecepta mea ut non subia- 22:9 They shall therefore keep mine ordinance, lest they bear sin for it, and die therefore, if they profane it: I the LORD do sanctify them.

ceant peccato et moriantur in sanctuario cum polluerint illud ego Dominus qui sanctifico eos 22:10 omnis alienigena 22:10 There shall no stranger eat of the holy thing: a so-

journer of the priest, or an hired servant, shall not eat of the holy thing.

non comedet de sanctificatis inquilinus sacerdotis et mercennarius non uescentur ex eis 22:11 quem autem sacerdos 22:11 But if the priest buy any soul with his money, he

shall eat of it, and he that is born in his house: they shall eat of his meat.

emerit et qui uernaculus domus eius fuerit hii comedent ex eis 22:12 si filia sacerdotis cuilibet ex populo nupta fue- 22:12 If the priest’s daughter also be married unto a

stranger, she may not eat of an offering of the holy things.

rit de his quae sanctificata sunt et de primitiis non uescetur 22:13 sin autem uidua uel repudiata et absque liberis reuersa 22:13 But if the priest’s daughter be a widow, or divorced,

and have no child, and is returned unto her father’s house, as in her youth, she shall eat of her father’s meat: but there shall be no stranger eat thereof.

Biblia Sacra Vulgata

fuerit ad domum patris sui sicut puella consuerat aletur cibis patris sui omnis alienigena comedendi ex eis non habet po-

22:14 And if a man eat of the holy thing unwittingly, then testatem 22:14 qui comederit de sanctificatis per ignoran-

he shall put the fifth part thereof unto it, and shall give it unto the priest with the holy thing.

tiam addet quintam partem cum eo quod comedit et dabit

22:15 And they shall not profane the holy things of the sacerdoti in sanctuarium 22:15 nec contaminabunt sanctifi-

children of Israel, which they offer unto the LORD; 22:16 Or suffer them to bear the iniquity of trespass, when cata filiorum Israhel quae offerunt Domino 22:16 ne forte they eat their holy things: for I the LORD do sanctify them.

sustineant iniquitatem delicti sui cum sanctificata comede-

22:17 And the LORD spake unto Moses, saying, rint ego Dominus qui sanctifico eos 22:17 locutus est Domi-

22:18 Speak unto Aaron, and to his sons, and unto all the nus ad Mosen dicens 22:18 loquere ad Aaron et filios eius children of Israel, and say unto them, Whatsoever he be of the house of Israel, or of the strangers in Israel, that will offer his oblation for all his vows, and for all his freewill offerings, which they will offer unto the LORD for a burnt offering;

et ad omnes filios Israhel dicesque ad eos homo de domo Israhel et de aduenis qui habitant apud uos qui obtulerit oblationem suam uel uota soluens uel sponte offerens quicquid

22:19 Ye shall offer at your own will a male without blem- illud obtulerit in holocaustum Domini 22:19 ut offeratur per

ish, of the beeves, of the sheep, or of the goats.

uos masculus inmaculatus erit ex bubus et ex ouibus et ex

22:20 But whatsoever hath a blemish, that shall ye not of- capris 22:20 si maculam habuerit non offeretis neque erit

fer: for it shall not be acceptable for you.

22:21 And whosoever offereth a sacrifice of peace offer- acceptabile 22:21 homo qui obtulerit uictimam pacificorum

ings unto the LORD to accomplish his vow, or a freewill offering in beeves or sheep, it shall be perfect to be ac-

cepted; there shall be no blemish therein. Domino uel uota soluens uel sponte offerens tam de bubus

quam de ouibus inmaculatum offeret ut acceptabile sit om-

22:22 Blind, or broken, or maimed, or having a wen, nis macula non erit in eo 22:22 si caecum fuerit si fractum

or scurvy, or scabbed, ye shall not offer these unto the LORD, nor make an offering by fire of them upon the al-

tar unto the LORD. si cicatricem habens si papulas aut scabiem uel inpetiginem

non offeretis ea Domino neque adolebitis ex eis super al-

22:23 Either a bullock or a lamb that hath any thing super- tare Domini 22:23 bouem et ouem aure et cauda amputatis fluous or lacking in his parts, that mayest thou offer for a freewill offering; but for a vow it shall not be accepted.

uoluntarie offerre potes uotum autem ex his solui non po-

22:24 Ye shall not offer unto the LORD that which is test 22:24 omne animal quod uel contritis uel tunsis uel sec-

bruised, or crushed, or broken, or cut; neither shall ye make any offering thereof in your land.

tis ablatisque testiculis est non offeretis Domino et in terra

22:25 Neither from a stranger’s hand shall ye offer the uestra hoc omnino ne faciatis 22:25 de manu alienigenae

bread of your God of any of these; because their corruption is in them, and blemishes be in them: they shall not

be accepted for you. non offeretis panes Deo uestro et quicquid aliud dare uo-

luerint quia corrupta et maculata sunt omnia non suscipietis

22:26 And the LORD spake unto Moses, saying, ea 22:26 locutusque est Dominus ad Mosen dicens 22:27

22:27 When a bullock, or a sheep, or a goat, is brought forth, then it shall be seven days under the dam; and from the eighth day and thenceforth it shall be accepted for an offering made by fire unto the LORD.

bos ouis et capra cum genita fuerint septem diebus erunt sub ubere matris suae die autem octauo et deinceps offerri po-

22:28 And whether it be cow, or ewe, ye shall not kill it terunt Domino 22:28 siue illa bos siue ouis non immola-

and her young both in one day.

22:29 And when ye will offer a sacrifice of thanksgiving buntur una die cum fetibus suis 22:29 si immolaueritis ho-

unto the LORD, offer it at your own will.

Liber Leuiticus

stiam pro gratiarum actione Domino ut possit esse placabilis 22:30 eodem die comedetis eam non remanebit quic- 22:30 On the same day it shall be eaten up; ye shall leave

none of it until the morrow: I am the LORD.

quam in mane alterius diei ego Dominus 22:31 custodite 22:31 Therefore shall ye keep my commandments, and do

them: I am the LORD.

mandata mea et facite ea ego Dominus 22:32 ne polluatis 22:32 Neither shall ye profane my holy name; but I will

be hallowed among the children of Israel: I am the LORD which hallow you,

nomen meum sanctum ut sanctificer in medio filiorum Israhel ego Dominus qui sanctifico uos 22:33 et eduxi de terra 22:33 That brought you out of the land of Egypt, to be

your God: I am the LORD.

Aegypti ut essem uobis in Deum ego Dominus

23:1 locutus est Dominus ad Mosen dicens 23:2 loquere 23:1 And the LORD spake unto Moses, saying,

23:2 Speak unto the children of Israel, and say unto them, Concerning the feasts of the LORD, which ye shall proclaim to be holy convocations, even these are my feasts.

filiis Israhel et dices ad eos hae sunt feriae Domini quas uocabitis sanctas 23:3 sex diebus facietis opus dies septimus 23:3 Six days shall work be done: but the seventh day is

the sabbath of rest, an holy convocation; ye shall do no work therein: it is the sabbath of the LORD in all your

quia sabbati requies est uocabitur sanctus omne opus non dwellings. facietis in eo sabbatum Domini est in cunctis habitationibus uestris 23:4 hae sunt ergo feriae Domini sanctae quas 23:4 These are the feasts of the LORD, even holy convo-

cations, which ye shall proclaim in their seasons.

celebrare debetis temporibus suis 23:5 mense primo quar- 23:5 In the fourteenth day of the first month at even is the

LORD’s passover.

tadecima die mensis ad uesperum phase Domini est 23:6 et 23:6 And on the fifteenth day of the same month is the

feast of unleavened bread unto the LORD: seven days ye must eat unleavened bread.

quintadecima die mensis huius sollemnitas azymorum Domini est septem diebus azyma comedetis 23:7 dies primus 23:7 In the first day ye shall have an holy convocation: ye

shall do no servile work therein.

erit uobis celeberrimus sanctusque omne opus seruile non facietis in eo 23:8 sed offeretis sacrificium in igne Domino 23:8 But ye shall offer an offering made by fire unto the

LORD seven days: in the seventh day is an holy convocation: ye shall do no servile work therein.

septem diebus dies autem septimus erit celebrior et sanctior nullumque seruile opus fiet in eo 23:9 locutusque est Do- 23:9 And the LORD spake unto Moses, saying,

minus ad Mosen dicens 23:10 loquere filiis Israhel et di- 23:10 Speak unto the children of Israel, and say unto them, When ye be come into the land which I give unto you, and shall reap the harvest thereof, then ye shall bring a sheaf

ces ad eos cum ingressi fueritis terram quam ego dabo uo- of the firstfruits of your harvest unto the priest: bis et messueritis segetem feretis manipulos spicarum primitias messis uestrae ad sacerdotem 23:11 qui eleuabit fa- 23:11 And he shall wave the sheaf before the LORD, to

be accepted for you: on the morrow after the sabbath the priest shall wave it.

sciculum coram Domino ut acceptabile sit pro uobis altero die sabbati et sanctificabit illum 23:12 atque in eodem die 23:12 And ye shall offer that day when ye wave the sheaf

an he lamb without blemish of the first year for a burnt offering unto the LORD.

quo manipulus consecratur caedetur agnus inmaculatus anniculus in holocaustum Domini 23:13 et libamenta offe- 23:13 And the meat offering thereof shall be two tenth

deals of fine flour mingled with oil, an offering made by fire unto the LORD for a sweet savour: and the drink of-

rentur cum eo duae decimae similae conspersae oleo in in- fering thereof shall be of wine, the fourth part of an hin. censum Domini odoremque suauissimum liba quoque uini quarta pars hin 23:14 panem et pulentam et pultes non co- 23:14 And ye shall eat neither bread, nor parched corn,

nor green ears, until the selfsame day that ye have brought an offering unto your God: it shall be a statute for ever

medetis ex segete usque ad diem qua offeratis ex ea Deo throughout your generations in all your dwellings.

Biblia Sacra Vulgata

uestro praeceptum est sempiternum in generationibus cunc-

23:15 And ye shall count unto you from the morrow after tisque habitaculis uestris 23:15 numerabitis ergo ab altero

the sabbath, from the day that ye brought the sheaf of the wave offering; seven sabbaths shall be complete:

die sabbati in quo obtulistis manipulum primitiarum septem

23:16 Even unto the morrow after the seventh sabbath ebdomadas plenas 23:16 usque ad alteram diem expletionis

shall ye number fifty days; and ye shall offer a new meat offering unto the LORD.

ebdomadae septimae id est quinquaginta dies et sic offeretis

23:17 Ye shall bring out of your habitations two wave sacrificium nouum Domino 23:17 ex omnibus habitaculis

loaves of two tenth deals; they shall be of fine flour; they shall be baken with leaven; they are the firstfruits unto the

LORD. uestris panes primitiarum duos de duabus decimis similae 23:18 And ye shall offer with the bread seven lambs with- fermentatae quos coquetis in primitias Domini 23:18 offe-

out blemish of the first year, and one young bullock, and two rams: they shall be for a burnt offering unto the LORD, with their meat offering, and their drink offerings, even an offering made by fire, of sweet savour unto the LORD.

retisque cum panibus septem agnos inmaculatos anniculos et uitulum de armento unum et arietes duos et erunt in holocausto cum libamentis suis in odorem suauissimum Do-

23:19 Then ye shall sacrifice one kid of the goats for a sin mino 23:19 facietis et hircum pro peccato duosque agnos offering, and two lambs of the first year for a sacrifice of peace offerings.

anniculos hostias pacificorum 23:20 cumque eleuauerit eos

23:20 And the priest shall wave them with the bread of the firstfruits for a wave offering before the LORD, with the two lambs: they shall be holy to the LORD for the priest.

sacerdos cum panibus primitiarum coram Domino cedent in

23:21 And ye shall proclaim on the selfsame day, that it usum eius 23:21 et uocabitis hunc diem celeberrimum at-

may be an holy convocation unto you: ye shall do no servile work therein: it shall be a statute for ever in all

your dwellings throughout your generations. que sanctissimum omne opus seruile non facietis in eo le-

gitimum sempiternum erit in cunctis habitaculis et genera-

23:22 And when ye reap the harvest of your land, thou tionibus uestris 23:22 postquam autem messueritis segetem

shalt not make clean riddance of the corners of thy field when thou reapest, neither shalt thou gather any gleaning of thy harvest: thou shalt leave them unto the poor, and to the stranger: I am the LORD your God.

terrae uestrae non secabitis eam usque ad solum nec remanentes spicas colligetis sed pauperibus et peregrinis dimitte-

23:23 And the LORD spake unto Moses, saying, tis eas ego Dominus Deus uester 23:23 locutusque est Do-

23:24 Speak unto the children of Israel, saying, In the sev- minus ad Mosen dicens 23:24 loquere filiis Israhel mense enth month, in the first day of the month, shall ye have a sabbath, a memorial of blowing of trumpets, an holy con-

vocation. septimo prima die mensis erit uobis sabbatum memorabile 23:25 Ye shall do no servile work therein: but ye shall clangentibus tubis et uocabitur sanctum 23:25 omne opus

offer an offering made by fire unto the LORD.

seruile non facietis in eo et offeretis holocaustum Domino

23:26 And the LORD spake unto Moses, saying, 23:26 locutusque est Dominus ad Mosen dicens 23:27 de-

23:27 Also on the tenth day of this seventh month there shall be a day of atonement: it shall be an holy convocation unto you; and ye shall afflict your souls, and offer an offering made by fire unto the LORD.

cimo die mensis huius septimi dies expiationum erit celeberrimus et uocabitur sanctus adfligetisque animas uestras

23:28 And ye shall do no work in that same day: for it is in eo et offeretis holocaustum Domino 23:28 omne opus

a day of atonement, to make an atonement for you before the LORD your God.

non facietis in tempore diei huius quia dies propitiationis est

23:29 For whatsoever soul it be that shall not be afflicted in ut propitietur uobis Dominus Deus uester 23:29 omnis anthat same day, he shall be cut off from among his people.

ima quae adflicta non fuerit die hoc peribit de populis suis

23:30 And whatsoever soul it be that doeth any work in 23:30 et quae operis quippiam fecerit delebo eam de populo

that same day, the same soul will I destroy from among his people.

Liber Leuiticus

suo 23:31 nihil ergo operis facietis in eo legitimum sem- 23:31 Ye shall do no manner of work: it shall be a statute

for ever throughout your generations in all your dwellings.

piternum erit uobis in cunctis generationibus et habitationibus uestris 23:32 sabbatum requietionis est adfligetis ani- 23:32 It shall be unto you a sabbath of rest, and ye shall

afflict your souls: in the ninth day of the month at even, from even unto even, shall ye celebrate your sabbath.

mas uestras die nono mensis a uespero usque ad uesperum celebrabitis sabbata uestra 23:33 et locutus est Dominus ad 23:33 And the LORD spake unto Moses, saying, Mosen dicens 23:34 loquere filiis Israhel a quintodecimo 23:34 Speak unto the children of Israel, saying, The fif-

teenth day of this seventh month shall be the feast of tabernacles for seven days unto the LORD.

die mensis huius septimi erunt feriae tabernaculorum septem diebus Domino 23:35 dies primus uocabitur celeberrimus 23:35 On the first day shall be an holy convocation: ye

shall do no servile work therein.

atque sanctissimus omne opus seruile non facietis 23:36 et 23:36 Seven days ye shall offer an offering made by fire

unto the LORD: on the eighth day shall be an holy convocation unto you; and ye shall offer an offering made by fire unto the LORD: it is a solemn assembly; and ye shall do no servile work therein.

septem diebus offeretis holocausta Domino dies quoque octauus erit celeberrimus atque sanctissimus et offeretis holocaustum Domino est enim coetus atque collectae omne opus seruile non facietis in eo 23:37 hae sunt feriae Domini quas 23:37 These are the feasts of the LORD, which ye shall

proclaim to be holy convocations, to offer an offering made by fire unto the LORD, a burnt offering, and a meat offering, a sacrifice, and drink offerings, every thing upon his day:

uocabitis celeberrimas et sanctissimas offeretisque in eis oblationes Domino holocausta et libamenta iuxta ritum uniuscuiusque diei 23:38 exceptis sabbatis Domini donisque ue- 23:38 Beside the sabbaths of the LORD, and beside your

gifts, and beside all your vows, and beside all your freewill offerings, which ye give unto the LORD.

stris et quae offertis ex uoto uel quae sponte tribuitis Domino 23:39 a quintodecimo ergo die mensis septimi quando con- 23:39 Also in the fifteenth day of the seventh month, when

ye have gathered in the fruit of the land, ye shall keep a feast unto the LORD seven days: on the first day shall be

gregaueritis omnes fructus terrae uestrae celebrabitis ferias a sabbath, and on the eighth day shall be a sabbath. Domini septem diebus die primo et die octauo erit sabbatum id est requies 23:40 sumetisque uobis die primo fructus 23:40 And ye shall take you on the first day the boughs

of goodly trees, branches of palm trees, and the boughs of thick trees, and willows of the brook; and ye shall rejoice

arboris pulcherrimae spatulasque palmarum et ramos ligni before the LORD your God seven days. densarum frondium et salices de torrente et laetabimini coram Domino Deo uestro 23:41 celebrabitisque sollemnita- 23:41 And ye shall keep it a feast unto the LORD seven

days in the year. It shall be a statute for ever in your generations: ye shall celebrate it in the seventh month.

tem eius septem diebus per annum legitimum sempiternum erit in generationibus uestris mense septimo festa celebrabitis 23:42 et habitabitis in umbraculis septem diebus omnis 23:42 Ye shall dwell in booths seven days; all that are Is-

raelites born shall dwell in booths:

qui de genere est Israhel manebit in tabernaculis 23:43 ut 23:43 That your generations may know that I made the

children of Israel to dwell in booths, when I brought them out of the land of Egypt: I am the LORD your God.

discant posteri uestri quod in tabernaculis habitare fecerim filios Israhel cum educerem eos de terra Aegypti ego Dominus Deus uester 23:44 locutusque est Moses super sollem- 23:44 And Moses declared unto the children of Israel the

feasts of the LORD.

nitatibus Domini ad filios Israhel

24:1 et locutus est Dominus ad Mosen dicens 24:2 prae- 24:1 And the LORD spake unto Moses, saying,

24:2 Command the children of Israel, that they bring unto thee pure oil olive beaten for the light, to cause the lamps to burn continually.

Biblia Sacra Vulgata

cipe filiis Israhel ut adferant tibi oleum de oliuis purissimum

24:3 Without the vail of the testimony, in the tabernacle ac lucidum ad concinnandas lucernas iugiter 24:3 extra ue-

of the congregation, shall Aaron order it from the evening unto the morning before the LORD continually: it shall be

a statute for ever in your generations. lum testimonii in tabernaculo foederis ponetque eas Aaron

a uespere usque in mane coram Domino cultu rituque per-

24:4 He shall order the lamps upon the pure candlestick petuo in generationibus uestris 24:4 super candelabro mun-

before the LORD continually.

24:5 And thou shalt take fine flour, and bake twelve cakes dissimo ponentur semper in conspectu Domini 24:5 acci-

thereof: two tenth deals shall be in one cake.

pies quoque similam et coques ex ea duodecim panes qui

24:6 And thou shalt set them in two rows, six on a row, singuli habebunt duas decimas 24:6 quorum senos altrinse-

upon the pure table before the LORD. 24:7 And thou shalt put pure frankincense upon each row, cus super mensam purissimam coram Domino statues 24:7 that it may be on the bread for a memorial, even an offering made by fire unto the LORD.

et pones super eos tus lucidissimum ut sit panis in monu-

24:8 Every sabbath he shall set it in order before the mentum oblationis Domini 24:8 per singula sabbata muta-

LORD continually, being taken from the children of Israel by an everlasting covenant.

buntur coram Domino suscepti a filiis Israhel foedere sem-

24:9 And it shall be Aaron’s and his sons’; and they shall piterno 24:9 eruntque Aaron et filiorum eius ut comedant eat it in the holy place: for it is most holy unto him of the offerings of the LORD made by fire by a perpetual statute.

eos in loco sancto quia sanctum sanctorum est de sacrificiis

24:10 And the son of an Israelitish woman, whose father Domini iure perpetuo 24:10 ecce autem egressus filius mu-

was an Egyptian, went out among the children of Israel: and this son of the Israelitish woman and a man of Israel

strove together in the camp; lieris israhelitis quem pepererat de uiro aegyptio inter filios 24:11 And the Israelitish woman’s son blasphemed the Israhel iurgatus est in castris cum uiro israhelite 24:11 cum-

name of the Lord, and cursed. And they brought him unto Moses: (and his mother’s name was Shelomith, the daugh-

ter of Dibri, of the tribe of Dan:) que blasphemasset nomen et maledixisset ei adductus est ad

Mosen uocabatur autem mater eius Salumith filia Dabri de

24:12 And they put him in ward, that the mind of the tribu Dan 24:12 miseruntque eum in carcerem donec nos-

LORD might be shewed them.

24:13 And the LORD spake unto Moses, saying, sent quid iuberet Dominus 24:13 qui locutus est ad Mosen 24:14 Bring forth him that hath cursed without the camp; 24:14 dicens educ blasphemum extra castra et ponant omnes

and let all that heard him lay their hands upon his head, and let all the congregation stone him.

qui audierunt manus suas super caput eius et lapidet eum

24:15 And thou shalt speak unto the children of Israel, populus uniuersus 24:15 et ad filios Israhel loqueris homo

saying, Whosoever curseth his God shall bear his sin.

24:16 And he that blasphemeth the name of the LORD, he qui maledixerit Deo suo portabit peccatum suum 24:16 et

shall surely be put to death, and all the congregation shall certainly stone him: as well the stranger, as he that is born in the land, when he blasphemeth the name of the Lord, shall be put to death.

qui blasphemauerit nomen Domini morte moriatur lapidibus opprimet eum omnis multitudo siue ille ciuis seu peregrinus fuerit qui blasphemauerit nomen Domini morte moria-

24:17 And he that killeth any man shall surely be put to tur 24:17 qui percusserit et occiderit hominem morte mo-

death.

24:18 And he that killeth a beast shall make it good; beast riatur 24:18 qui percusserit animal reddat uicarium id est

for beast.

24:19 And if a man cause a blemish in his neighbour; as animam pro anima 24:19 qui inrogauerit maculam cuilibet

he hath done, so shall it be done to him;

24:20 Breach for breach, eye for eye, tooth for tooth: as ciuium suorum sicut fecit fiet ei 24:20 fracturam pro frac-

he hath caused a blemish in a man, so shall it be done to him again.

tura oculum pro oculo dentem pro dente restituet qualem in-

Liber Leuiticus

flixerit maculam talem sustinere cogetur 24:21 qui percus- 24:21 And he that killeth a beast, he shall restore it: and

he that killeth a man, he shall be put to death.

serit iumentum reddet aliud qui percusserit hominem punietur 24:22 aequum iudicium sit inter uos siue peregrinus siue 24:22 Ye shall have one manner of law, as well for the

stranger, as for one of your own country: for I am the LORD your God.

ciuis peccauerit quia ego sum Dominus Deus uester 24:23

24:23 And Moses spake to the children of Israel, that they should bring forth him that had cursed out of the camp, and stone him with stones. And the children of Israel did as the LORD commanded Moses.

locutusque est Moses ad filios Israhel et eduxerunt eum qui blasphemauerat extra castra ac lapidibus oppresserunt feceruntque filii Israhel sicut praeceperat Dominus Mosi

25:1 locutusque est Dominus ad Mosen in monte Sinai di- 25:1 And the LORD spake unto Moses in mount Sinai,

saying,

cens 25:2 loquere filiis Israhel et dices ad eos quando in- 25:2 Speak unto the children of Israel, and say unto them,

When ye come into the land which I give you, then shall the land keep a sabbath unto the LORD.

gressi fueritis terram quam ego dabo uobis sabbatizet sabba-

tum Domini 25:3 sex annis seres agrum tuum et sex annis 25:3 Six years thou shalt sow thy field, and six years thou

shalt prune thy vineyard, and gather in the fruit thereof;

putabis uineam tuam colligesque fructus eius 25:4 septimo 25:4 But in the seventh year shall be a sabbath of rest unto

the land, a sabbath for the LORD: thou shalt neither sow thy field, nor prune thy vineyard.

autem anno sabbatum erit terrae requietionis Domini agrum non seres et uineam non putabis 25:5 quae sponte gignit 25:5 That which groweth of its own accord of thy harvest

thou shalt not reap, neither gather the grapes of thy vine undressed: for it is a year of rest unto the land.

humus non metes et uuas primitiarum tuarum non colliges quasi uindemiam annus enim requietionis terrae est 25:6 25:6 And the sabbath of the land shall be meat for you;

for thee, and for thy servant, and for thy maid, and for thy hired servant, and for thy stranger that sojourneth with

sed erunt uobis in cibum tibi et seruo tuo ancillae et mer- thee. cennario tuo et aduenae qui peregrinantur apud te 25:7 iu- 25:7 And for thy cattle, and for the beast that are in thy

land, shall all the increase thereof be meat.

mentis tuis et pecoribus omnia quae nascuntur praebebunt cibum 25:8 numerabis quoque tibi septem ebdomades an- 25:8 And thou shalt number seven sabbaths of years unto

thee, seven times seven years; and the space of the seven sabbaths of years shall be unto thee forty and nine years.

norum id est septem septies quae simul faciunt annos quadraginta nouem 25:9 et clanges bucina mense septimo de- 25:9 Then shalt thou cause the trumpet of the jubile to

sound on the tenth day of the seventh month, in the day of atonement shall ye make the trumpet sound throughout all

cima die mensis propitiationis tempore in uniuersa terra ues- your land. tra 25:10 sanctificabisque annum quinquagesimum et uo- 25:10 And ye shall hallow the fiftieth year, and proclaim

liberty throughout all the land unto all the inhabitants thereof: it shall be a jubile unto you; and ye shall return every man unto his possession, and ye shall return every man unto his family.

cabis remissionem cunctis habitatoribus terrae tuae ipse est enim iobeleus reuertetur homo ad possessionem suam et unusquisque rediit ad familiam pristinam 25:11 quia iobe- 25:11 A jubile shall that fiftieth year be unto you: ye shall

not sow, neither reap that which groweth of itself in it, nor gather the grapes in it of thy vine undressed.

leus est et quinquagesimus annus non seretis neque metetis sponte in agro nascentia et primitias uindemiae non colligetis 25:12 ob sanctificationem iobelei sed statim ablata co- 25:12 For it is the jubile; it shall be holy unto you: ye shall

eat the increase thereof out of the field.

medetis 25:13 anno iobelei rediint omnes ad possessiones 25:13 In the year of this jubile ye shall return every man

unto his possession.

suas 25:14 quando uendes quippiam ciui tuo uel emes ab eo 25:14 And if thou sell ought unto thy neighbour, or buyest ought of thy neighbour’s hand, ye shall not oppress one another:

ne contristes fratrem tuum sed iuxta numerum annorum io-

Biblia Sacra Vulgata

25:15 According to the number of years after the jubile belei emes ab eo 25:15 et iuxta supputationem frugum uen-

thou shalt buy of thy neighbour, and according unto the number of years of the fruits he shall sell unto thee:

det tibi 25:16 quanto plus anni remanserint post iobeleum

25:16 According to the multitude of years thou shalt increase the price thereof, and according to the fewness of years thou shalt diminish the price of it: for according to the number of the years of the fruits doth he sell unto thee.

tanto crescet et pretium et quanto minus temporis numeraueris tanto minoris et emptio constabit tempus enim frugum

25:17 Ye shall not therefore oppress one another; but thou uendet tibi 25:17 nolite adfligere contribules uestros sed ti-

shalt fear thy God:for I am the LORD your God.

meat unusquisque Deum suum quia ego Dominus Deus ue-

25:18 Wherefore ye shall do my statutes, and keep my ster 25:18 facite praecepta mea et iudicia custodite et imp-

judgments, and do them; and ye shall dwell in the land in safety.

lete ea ut habitare possitis in terra absque ullo pauore 25:19

25:19 And the land shall yield her fruit, and ye shall eat your fill, and dwell therein in safety.

et gignat uobis humus fructus suos quibus uescamini usque

25:20 And if ye shall say, What shall we eat the seventh ad saturitatem nullius impetum formidantes 25:20 quod si

year? behold, we shall not sow, nor gather in our increase:

dixeritis quid comedemus anno septimo si non seruerimus

25:21 Then I will command my blessing upon you in the neque collegerimus fruges nostras 25:21 dabo benedictio-

sixth year, and it shall bring forth fruit for three years.

nem meam uobis anno sexto et faciet fructus trium anno-

25:22 And ye shall sow the eighth year, and eat yet of old rum 25:22 seretisque anno octauo et comedetis ueteres fru-

fruit until the ninth year; until her fruits come in ye shall eat of the old store.

ges usque ad nonum annum donec noua nascantur edetis ue-

25:23 The land shall not be sold for ever: for the land is tera 25:23 terra quoque non ueniet in perpetuum quia mea

mine, for ye are strangers and sojourners with me. 25:24 And in all the land of your possession ye shall grant est et uos aduenae et coloni mei estis 25:24 unde cuncta a redemption for the land.

regio possessionis uestrae sub redeeptionis condicione uen-

25:25 If thy brother be waxen poor, and hath sold away detur 25:25 si adtenuatus frater tuus uendiderit possessiun-

some of his possession, and if any of his kin come to redeem it, then shall he redeem that which his brother sold.

culam suam et uoluerit propinquus eius potest rediiere quod

25:26 And if the man have none to redeem it, and himself ille uendiderat 25:26 sin autem non habuerit proximum et

be able to redeem it;

25:27 Then let him count the years of the sale thereof, and ipse pretium ad rediiendum potuerit inuenire 25:27 conpu-

restore the overplus unto the man to whom he sold it; that he may return unto his possession.

tabuntur fructus ex eo tempore quo uendidit et quod reliquum est reddet emptori sicque recipiet possessionem suam

25:28 But if he be not able to restore it to him, then that 25:28 quod si non inuenerit manus eius ut reddat pretium

which is sold shall remain in the hand of him that hath bought it until the year of jubile: and in the jubile it shall

go out, and he shall return unto his possession. habebit emptor quod emerat usque ad annum iobeleum in

ipso enim omnis uenditio redii ad dominum et ad possesso-

25:29 And if a man sell a dwelling house in a walled city, rem pristinum 25:29 qui uendiderit domum intra urbis mu-

then he may redeem it within a whole year after it is sold; within a full year may he redeem it.

ros habebit licentiam rediiendi donec unus impleatur annus

25:30 And if it be not redeemed within the space of a full 25:30 si non redeeerit et anni circulus fuerit euolutus emp-

year, then the house that is in the walled city shall be established for ever to him that bought it throughout his gen-

erations: it shall not go out in the jubile. tor possidebit eam et posteri eius in perpetuum et rediii non 25:31 But the houses of the villages which have no wall poterit etiam in iobeleo 25:31 sin autem in uilla fuerit do-

round about them shall be counted as the fields of the country: they may be redeemed, and they shall go out in

the jubile. mus quae muros non habet agrorum iure uendetur si ante

Liber Leuiticus

redeepta non fuerit in iobeleo reuertetur ad dominum 25:32 25:32 Notwithstanding the cities of the Levites, and the

houses of the cities of their possession, may the Levites redeem at any time.

aedes Leuitarum quae in urbibus sunt semper possunt rediii 25:33 si redeeptae non fuerint in iobeleo reuertentur ad 25:33 And if a man purchase of the Levites, then the house

that was sold, and the city of his possession, shall go out in the year of jubile: for the houses of the cities of the Levites

dominos quia domus urbium leuiticarum pro possessionibus are their possession among the children of Israel. sunt inter filios Israhel 25:34 suburbana autem eorum non 25:34 But the field of the suburbs of their cities may not

be sold; for it is their perpetual possession.

uenient quia possessio sempiterna est 25:35 si adtenuatus 25:35 And if thy brother be waxen poor, and fallen in de-

cay with thee; then thou shalt relieve him: yea, though he be a stranger, or a sojourner; that he may live with thee.

fuerit frater tuus et infirmus manu et susceperis eum quasi aduenam et peregrinum et uixerit tecum 25:36 ne accipias 25:36 Take thou no usury of him, or increase: but fear thy

God; that thy brother may live with thee.

usuras ab eo nec amplius quam dedisti time Deum tuum ut uiuere possit frater tuus apud te 25:37 pecuniam tuam 25:37 Thou shalt not give him thy money upon usury, nor

lend him thy victuals for increase.

non dabis ei ad usuram et frugum superabundantiam non exiges 25:38 ego Dominus Deus uester qui eduxi uos de 25:38 I am the LORD your God, which brought you forth

out of the land of Egypt, to give you the land of Canaan, and to be your God.

terra Aegypti ut darem uobis terram Chanaan et essem uester Deus 25:39 si paupertate conpulsus uendiderit se tibi 25:39 And if thy brother that dwelleth by thee be waxen

poor, and be sold unto thee; thou shalt not compel him to serve as a bondservant:

frater tuus non eum opprimes seruitute famulorum 25:40

25:40 But as an hired servant, and as a sojourner, he shall be with thee, and shall serve thee unto the year of jubile.

sed quasi mercennarius et colonus erit usque ad annum iobeleum operabitur apud te 25:41 et postea egredietur cum 25:41 And then shall he depart from thee, both he and his

children with him, and shall return unto his own family, and unto the possession of his fathers shall he return.

liberis suis et reuertetur ad cognationem et ad possessionem

patrum suorum 25:42 mei enim serui sunt et ego eduxi eos 25:42 For they are my servants, which I brought forth out

of the land of Egypt: they shall not be sold as bondmen.

de terra Aegypti non uenient condicione seruorum 25:43 ne 25:43 Thou shalt not rule over him with rigour; but shalt

fear thy God.

adfligas eum per potentiam sed metuito Deum tuum 25:44 25:44 Both thy bondmen, and thy bondmaids, which thou

shalt have, shall be of the heathen that are round about you; of them shall ye buy bondmen and bondmaids.

seruus et ancilla sint uobis de nationibus quae in circuitu uestro sunt 25:45 et de aduenis qui peregrinantur apud uos uel 25:45 Moreover of the children of the strangers that do so-

journ among you, of them shall ye buy, and of their families that are with you, which they begat in your land: and

qui ex his nati fuerint in terra uestra hos habebitis famulos they shall be your possession. 25:46 et hereditario iure transmittetis ad posteros ac posside- 25:46 And ye shall take them as an inheritance for your

children after you, to inherit them for a possession; they shall be your bondmen for ever: but over your brethren the children of Israel, ye shall not rule one over another with rigour.

bitis in aeternum fratres autem uestros filios Israhel ne opprimatis per potentiam 25:47 si inualuerit apud uos manus 25:47 And if a sojourner or stranger wax rich by thee, and

thy brother that dwelleth by him wax poor, and sell himself unto the stranger or sojourner by thee, or to the stock of the

aduenae atque peregrini et adtenuatus frater tuus uendide- stranger’s family: rit se ei aut cuiquam de stirpe eius 25:48 post uenditionem 25:48 After that he is sold he may be redeemed again; one

of his brethren may redeem him:

potest rediii qui uoluerit ex fratribus suis rediiet eum 25:49 25:49 Either his uncle, or his uncle’s son, may redeem

him, or any that is nigh of kin unto him of his family may redeem him; or if he be able, he may redeem himself.

et patruus et patruelis et consanguineus et adfinis sin autem et ipse potuerit rediiet se 25:50 supputatis dumtaxat annis a 25:50 And he shall reckon with him that bought him from

the year that he was sold to him unto the year of jubile: and the price of his sale shall be according unto the number of years, according to the time of an hired servant shall it be with him.

tempore uenditionis suae usque ad annum iobeleum et pecu-

Biblia Sacra Vulgata

nia qua uenditus fuerat iuxta annorum numerum et rationem

25:51 If there be yet many years behind, according unto mercennarii supputata 25:51 si plures fuerint anni qui re-

them he shall give again the price of his redemption out of the money that he was bought for.

manent usque ad iobeleum secundum hos reddet et pretium

25:52 And if there remain but few years unto the year of 25:52 si pauci ponet rationem cum eo iuxta annorum nume-

jubile, then he shall count with him, and according unto his years shall he give him again the price of his redemp-

tion. rum et reddet emptori quod reliquum est annorum 25:53

25:53 And as a yearly hired servant shall he be with him: and the other shall not rule with rigour over him in thy sight.

quibus ante seruiuit mercedibus inputatis non adfliget eum

25:54 And if he be not redeemed in these years, then he uiolenter in conspectu tuo 25:54 quod si per haec rediii non

shall go out in the year of jubile, both he, and his children with him.

potuerit anno iobeleo egredietur cum liberis suis 25:55 mei

25:55 For unto me the children of Israel are servants; they are my servants whom I brought forth out of the land of Egypt: I am the LORD your God.

sunt enim serui filii Israhel quos eduxi de terra Aegypti

26:1 Ye shall make you no idols nor graven image, neither 26:1 ego Dominus Deus uester non facietis uobis idolum rear you up a standing image, neither shall ye set up any image of stone in your land, to bow down unto it: for I am

the LORD your God. et sculptile nec titulos erigetis nec insignem lapidem pone-

tis in terra uestra ut adoretis eum ego enim sum Dominus

26:2 Ye shall keep my sabbaths, and reverence my sanctu- Deus uester 26:2 custodite sabbata mea et pauete ad sanc-

ary: I am the LORD.

26:3 If ye walk in my statutes, and keep my command- tuarium meum ego Dominus 26:3 si in praeceptis meis am-

ments, and do them;

bulaueritis et mandata mea custodieritis et feceritis ea dabo

26:4 Then I will give you rain in due season, and the land uobis pluuias temporibus suis 26:4 et terra gignet germen

shall yield her increase, and the trees of the field shall yield their fruit.

suum et pomis arbores replebuntur 26:5 adprehendet mes-

26:5 And your threshing shall reach unto the vintage, and the vintage shall reach unto the sowing time: and ye shall eat your bread to the full, and dwell in your land safely.

sium tritura uindemiam et uindemia occupabit sementem et comedetis panem uestrum in saturitatem et absque pauore

26:6 And I will give peace in the land, and ye shall lie habitabitis in terra uestra 26:6 dabo pacem in finibus ue-

down, and none shall make you afraid: and I will rid evil beasts out of the land, neither shall the sword go through

your land. stris dormietis et non erit qui exterreat auferam malas be-26:7 And ye shall chase your enemies, and they shall fall stias et gladius non transibit terminos uestros 26:7 perse-

before you by the sword.

26:8 And five of you shall chase an hundred, and an hun- quemini inimicos uestros et corruent coram uobis 26:8 per-

dred of you shall put ten thousand to flight: and your enemies shall fall before you by the sword.

sequentur quinque de uestris centum alienos et centum ex uobis decem milia cadent inimici uestri in conspectu uestro

26:9 For I will have respect unto you, and make you fruit- gladio 26:9 respiciam uos et crescere faciam multiplicabi-

ful, and multiply you, and establish my covenant with you. 26:10 And ye shall eat old store, and bring forth the old mini et firmabo pactum meum uobiscum 26:10 comedetis because of the new.

uetustissima ueterum et uetera nouis superuenientibus pro-

26:11 And I set my tabernacle among you: and my soul icietis 26:11 ponam tabernaculum meum in medio uestri

shall not abhor you.

26:12 And I will walk among you, and will be your God, et non abiciet uos anima mea 26:12 ambulabo inter uos et

and ye shall be my people.

26:13 I am the LORD your God, which brought you forth ero uester Deus uosque eritis populus meus 26:13 ego Do-

out of the land of Egypt, that ye should not be their bondmen; and I have broken the bands of your yoke, and made

you go upright. minus Deus uester qui eduxi uos de terra Aegyptiorum ne seruiretis eis et qui confregi catenas ceruicum uestrarum ut

Liber Leuiticus

incederetis erecti 26:14 quod si non audieritis me nec fe- 26:14 But if ye will not hearken unto me, and will not do

all these commandments;

ceritis omnia mandata mea 26:15 si spreueritis leges meas 26:15 And if ye shall despise my statutes, or if your soul

abhor my judgments, so that ye will not do all my commandments, but that ye break my covenant:

et iudicia mea contempseritis ut non faciatis ea quae a me constituta sunt et ad irritum perducatis pactum meum 26:16 26:16 I also will do this unto you; I will even appoint over

you terror, consumption, and the burning ague, that shall consume the eyes, and cause sorrow of heart: and ye shall

ego quoque haec faciam uobis uisitabo uos uelociter in ege- sow your seed in vain, for your enemies shall eat it. state et ardore qui conficiat oculos uestros et consumat animas frustra seretis sementem quae ab hostibus deuorabitur 26:17 ponam faciem meam contra uos et corruetis coram ho- 26:17 And I will set my face against you, and ye shall be

slain before your enemies: they that hate you shall reign over you; and ye shall flee when none pursueth you.

stibus uestris et subiciemini his qui oderunt uos fugietis nemine persequente 26:18 sin autem nec sic oboedieritis mihi 26:18 And if ye will not yet for all this hearken unto me,

then I will punish you seven times more for your sins.

addam correptiones uestras septuplum propter peccata uestra 26:19 et conteram superbiam duritiae uestrae daboque 26:19 And I will break the pride of your power; and I will

make your heaven as iron, and your earth as brass:

caelum uobis desuper sicut ferrum et terram aeneam 26:20 26:20 And your strength shall be spent in vain: for your

land shall not yield her increase, neither shall the trees of the land yield their fruits.

consumetur in cassum labor uester non proferet terra germen nec arbores poma praebebunt 26:21 si ambulaueritis ex ad- 26:21 And if ye walk contrary unto me, and will not hear-

ken unto me; I will bring seven times more plagues upon you according to your sins.

uerso mihi nec uolueritis audire me addam plagas uestras usque in septuplum propter peccata uestra 26:22 emittamque 26:22 I will also send wild beasts among you, which shall

rob you of your children, and destroy your cattle, and make you few in number; and your high ways shall be

in uos bestias agri quae consumant et uos et pecora uestra desolate. et ad paucitatem cuncta rediiant desertaeque fiant uiae uestrae 26:23 quod si nec sic uolueritis recipere disciplinam 26:23 And if ye will not be reformed by me by these

things, but will walk contrary unto me;

sed ambulaueritis ex aduerso mihi 26:24 ego quoque contra 26:24 Then will I also walk contrary unto you, and will

punish you yet seven times for your sins.

uos aduersus incedam et percutiam uos septies propter peccata uestra 26:25 inducamque super uos gladium ultorem 26:25 And I will bring a sword upon you, that shall avenge

the quarrel of my covenant: and when ye are gathered together within your cities, I will send the pestilence among

foederis mei cumque confugeritis in urbes mittam pestilen- you; and ye shall be delivered into the hand of the enemy. tiam in medio uestri et trademini hostium manibus 26:26 26:26 And when I have broken the staff of your bread, ten

women shall bake your bread in one oven, and they shall deliver you your bread again by weight: and ye shall eat,

postquam confregero baculum panis uestri ita ut decem mu- and not be satisfied. lieres in uno clibano coquant panes et reddant eos ad pondus et comedetis et non saturabimini 26:27 sin autem nec 26:27 And if ye will not for all this hearken unto me, but

walk contrary unto me;

per haec audieritis me sed ambulaueritis contra me 26:28 et 26:28 Then I will walk contrary unto you also in fury; and

I, even I, will chastise you seven times for your sins.

ego incedam aduersum uos in furore contrario et corripiam uos septem plagis propter peccata uestra 26:29 ita ut come- 26:29 And ye shall eat the flesh of your sons, and the flesh

of your daughters shall ye eat.

datis carnes filiorum et filiarum uestrarum 26:30 destruam 26:30 And I will destroy your high places, and cut down

your images, and cast your carcases upon the carcases of your idols, and my soul shall abhor you.

excelsa uestra et simulacra confringam cadetis inter ruinas

Biblia Sacra Vulgata

26:31 And I will make your cities waste, and bring your idolorum uestrorum et abominabitur uos anima mea 26:31

sanctuaries unto desolation, and I will not smell the savour of your sweet odours.

in tantum ut urbes uestras rediiam in solitudinem et deserta faciam sanctuaria uestra nec recipiam ultra odorem suauissi-

26:32 And I will bring the land into desolation: and your mum 26:32 disperdamque terram uestram et stupebunt suenemies which dwell therein shall be astonished at it.

26:33 And I will scatter you among the heathen, and will per ea inimici uestri cum habitatores illius fuerint 26:33 uos

draw out a sword after you: and your land shall be desolate, and your cities waste.

autem dispergam in gentes et euaginabo post uos gladium

26:34 Then shall the land enjoy her sabbaths, as long as it eritque terra uestra deserta et ciuitates dirutae 26:34 tunc

lieth desolate, and ye be in your enemies’ land; even then shall the land rest, and enjoy her sabbaths.

placebunt terrae sabbata sua cunctis diebus solitudinis suae

26:35 As long as it lieth desolate it shall rest; because it quando fueritis 26:35 in terra hostili sabbatizabit et requie-

did not rest in your sabbaths, when ye dwelt upon it.

scet in sabbatis solitudinis suae eo quod non requieuerit in

26:36 And upon them that are left alive of you I will send sabbatis uestris quando habitabatis in ea 26:36 et qui de uo-

a faintness into their hearts in the lands of their enemies; and the sound of a shaken leaf shall chase them; and they shall flee, as fleeing from a sword; and they shall fall when none pursueth.

bis remanserint dabo pauorem in cordibus eorum in regionibus hostium terrebit eos sonitus folii uolantis et ita fugient

26:37 And they shall fall one upon another, as it were be- quasi gladium cadent nullo sequente 26:37 et corruent sin-

fore a sword, when none pursueth: and ye shall have no power to stand before your enemies.

guli super fratres suos quasi bella fugientes nemo uestrum

26:38 And ye shall perish among the heathen, and the land inimicis audebit resistere 26:38 peribitis inter gentes et ho-

of your enemies shall eat you up.

26:39 And they that are left of you shall pine away in their stilis uos terra consumet 26:39 quod si et de his aliqui re-

iniquity in your enemies’ lands; and also in the iniquities of their fathers shall they pine away with them.

manserint tabescent in iniquitatibus suis in terra inimicorum suorum et propter peccata patrum suorum et sua adfligentur

26:40 If they shall confess their iniquity, and the iniquity 26:40 donec confiteantur iniquitates suas et maiorum suorum of their fathers, with their trespass which they trespassed against me, and that also they have walked contrary unto

me; quibus praeuaricati sunt in me et ambulauerunt ex aduerso 26:41 And that I also have walked contrary unto them, mihi 26:41 ambulabo igitur et ego contra eos et inducam

and have brought them into the land of their enemies; if then their uncircumcised hearts be humbled, and they then

accept of the punishment of their iniquity: illos in terram hostilem donec erubescat incircumcisa mens 26:42 Then will I remember my covenant with Jacob, and eorum tunc orabunt pro impietatibus suis 26:42 et recorda-

also my covenant with Isaac, and also my covenant with Abraham will I remember; and I will remember the land.

bor foederis mei quod pepigi cum Iacob et Isaac et Abraham

26:43 The land also shall be left of them, and shall en- terrae quoque memor ero 26:43 quae cum relicta fuerit ab

joy her sabbaths, while she lieth desolate without them: and they shall accept of the punishment of their iniquity: because, even because they despised my judgments, and because their soul abhorred my statutes.

eis conplacebit sibi in sabbatis suis patiens solitudinem propter illos ipsi uero rogabunt pro peccatis suis eo quod abie-

26:44 And yet for all that, when they be in the land of their cerint iudicia mea et leges meas despexerint 26:44 et tamen

enemies, I will not cast them away, neither will I abhor them, to destroy them utterly, and to break my covenant

with them: for I am the LORD their God. etiam cum essent in terra hostili non penitus abieci eos ne-

que sic despexi ut consumerentur et irritum facerem pactum

26:45 But I will for their sakes remember the covenant meum cum eis ego enim sum Dominus Deus eorum 26:45 of their ancestors, whom I brought forth out of the land of Egypt in the sight of the heathen, that I might be their God: I am the LORD. These are the statutes and judgments and laws, which the LORD made between him and the children of Israel in mount Sinai by the hand of Moses.

et recordabor foederis mei pristini quando eduxi eos de terra

Liber Leuiticus

Aegypti in conspectu gentium ut essem Deus eorum ego Dominus Deus haec sunt praecepta atque iudicia et leges quas dedit Dominus inter se et inter filios Israhel in monte Sinai per manum Mosi

27:1 locutusque est Dominus ad Mosen dicens 27:2 lo- 27:1 And the LORD spake unto Moses, saying,

27:2 Speak unto the children of Israel, and say unto them, When a man shall make a singular vow, the persons shall be for the LORD by thy estimation.

quere filiis Israhel et dices ad eos homo qui uotum fecerit et spoponderit Deo animam suam sub aestimatione dabit pretium 27:3 si fuerit masculus a uicesimo usque ad sexagesi- 27:3 And thy estimation shall be of the male from twenty

years old even unto sixty years old, even thy estimation shall be fifty shekels of silver, after the shekel of the sanc-

mum annum dabit quinquaginta siclos argenti ad mensuram tuary. sanctuarii 27:4 si mulier triginta 27:5 a quinto autem anno 27:4 And if it be a female, then thy estimation shall be

thirty shekels.

27:5 And if it be from five years old even unto twenty years old, then thy estimation shall be of the male twenty shekels, and for the female ten shekels.

usque ad uicesimum masculus dabit uiginti siclos femina decem 27:6 ab uno mense usque ad annum quintum pro ma- _{27:6 And if it be from a month old even unto five years}

old, then thy estimation shall be of the male five shekels of silver, and for the female thy estimation shall be three

sculo dabuntur quinque sicli pro femina tres 27:7 sexage- _{shekels of silver.}

27:7 And if it be from sixty years old and above; if it be a male, then thy estimation shall be fifteen shekels, and for the female ten shekels.

narius et ultra masculus dabit quindecim siclos femina decem 27:8 si pauper fuerit et aestimationem reddere non ua- 27:8 But if he be poorer than thy estimation, then he shall

present himself before the priest, and the priest shall value him; according to his ability that vowed shall the priest

lebit stabit coram sacerdote et quantum ille aestimauerit et value him. uiderit eum posse reddere tantum dabit 27:9 animal autem 27:9 And if it be a beast, whereof men bring an offering

unto the LORD, all that any man giveth of such unto the LORD shall be holy.

quod immolari potest Domino si quis uouerit sanctum erit 27:10 et mutari non poterit id est nec melius malo nec peius 27:10 He shall not alter it, nor change it, a good for a bad,

or a bad for a good: and if he shall at all change beast for beast, then it and the exchange thereof shall be holy.

bono quod si mutauerit et ipsum quod mutatum est et illud pro quo mutatum est consecratum erit Domino 27:11 ani- 27:11 And if it be any unclean beast, of which they do not

offer a sacrifice unto the LORD, then he shall present the beast before the priest:

mal inmundum quod immolari Domino non potest si quis uouerit adducetur ante sacerdotem 27:12 qui diiudicans ut- 27:12 And the priest shall value it, whether it be good or

bad: as thou valuest it, who art the priest, so shall it be.

rum bonum an malum sit statuet pretium 27:13 quod si dare 27:13 But if he will at all redeem it, then he shall add a

fifth part thereof unto thy estimation.

uoluerit is qui offert addet supra aestimationis quintam partem 27:14 homo si uouerit domum suam et sanctificauerit 27:14 And when a man shall sanctify his house to be holy

unto the LORD, then the priest shall estimate it, whether it be good or bad: as the priest shall estimate it, so shall it

Domino considerabit eam sacerdos utrum bona an mala sit stand. et iuxta pretium quod ab eo fuerit constitutum uenundabitur 27:15 sin autem ille qui uouerat uoluerit rediiere eam 27:15 And if he that sanctified it will redeem his house,

then he shall add the fifth part of the money of thy estimation unto it, and it shall be his.

dabit quintam partem aestimationis supra et habebit domum 27:16 quod si agrum possessionis suae uouerit et consecraue- 27:16 And if a man shall sanctify unto the LORD some

part of a field of his possession, then thy estimation shall be according to the seed thereof: an homer of barley seed

rit Domino iuxta mensuram sementis aestimabitur pretium si shall be valued at fifty shekels of silver. triginta modiis hordei seritur terra quinquaginta siclis ueniet

Biblia Sacra Vulgata

27:17 If he sanctify his field from the year of jubile, ac- argenti 27:17 si statim ab anno incipientis iobelei uouerit

cording to thy estimation it shall stand.

27:18 But if he sanctify his field after the jubile, then the agrum quanto ualere potest tanto aestimabitur 27:18 sin au-

priest shall reckon unto him the money according to the years that remain, even unto the year of the jubile, and it

shall be abated from thy estimation. tem post aliquantum temporis supputabit sacerdos pecuniam

iuxta annorum qui reliqui sunt numerum usque ad iobeleum

27:19 And if he that sanctified the field will in any wise et detrahetur ex pretio 27:19 quod si uoluerit rediiere agrum

redeem it, then he shall add the fifth part of the money of thy estimation unto it, and it shall be assured to him.

ille qui uouerat addet quintam partem aestimatae pecuniae

27:20 And if he will not redeem the field, or if he have et possidebit eum 27:20 sin autem noluerit rediiere sed al-

sold the field to another man, it shall not be redeemed any more.

teri cuilibet fuerit uenundatus ultra eum qui uouerat rediiere

27:21 But the field, when it goeth out in the jubile, shall non poterit 27:21 quia cum iobelei uenerit dies sanctificatus

be holy unto the LORD, as a field devoted; the possession thereof shall be the priest’s.

erit Domino et possessio consecrata ad ius pertinet sacerdo-

27:22 And if a man sanctify unto the LORD a field which tum 27:22 si ager emptus et non de possessione maiorum

he hath bought, which is not of the fields of his possession;

27:23 Then the priest shall reckon unto him the worth of sanctificatus fuerit Domino 27:23 supputabit sacerdos iuxta

thy estimation, even unto the year of the jubile: and he shall give thine estimation in that day, as a holy thing unto

the LORD. annorum numerum usque ad iobeleum pretium et dabit ille 27:24 In the year of the jubile the field shall return unto qui uouerat eum Domino 27:24 in iobeleo autem reuertetur

him of whom it was bought, even to him to whom the possession of the land did belong.

ad priorem dominum qui uendiderat eum et habuerat in sor-

27:25 And all thy estimations shall be according to the tem possessionis suae 27:25 omnis aestimatio siclo sanc-

shekel of the sanctuary: twenty gerahs shall be the shekel.

27:26 Only the firstling of the beasts, which should be the tuarii ponderabitur siclus uiginti obolos habet 27:26 primo-

LORD’s firstling, no man shall sanctify it; whether it be ox, or sheep: it is the LORD’s.

genita quae ad Dominum pertinent nemo sanctificare poterit

27:27 And if it be of an unclean beast, then he shall redeem et uouere siue bos siue ouis fuerit Domini sunt 27:27 quod it according to thine estimation, and shall add a fifth part of it thereto: or if it be not redeemed, then it shall be sold

according to thy estimation. si inmundum est animal rediiet qui obtulit iuxta aestimatio-

nem tuam et addet quintam partem pretii si rediiere noluerit

27:28 Notwithstanding no devoted thing, that a man shall uendetur alteri quantocumque a te fuerit aestimatum 27:28

devote unto the LORD of all that he hath, both of man and beast, and of the field of his possession, shall be sold or redeemed: every devoted thing is most holy unto the LORD.

omne quod Domino consecratur siue homo fuerit siue animal siue ager non ueniet nec rediii poterit quicquid semel

27:29 None devoted, which shall be devoted of men, shall fuerit consecratum sanctum sanctorum erit Domino 27:29

be redeemed; but shall surely be put to death.

et omnis consecratio quae offertur ab homine non rediietur

27:30 And all the tithe of the land, whether of the seed of sed morte morietur 27:30 omnes decimae terrae siue de fru-

the land, or of the fruit of the tree, is the LORD’s: it is holy unto the LORD.

gibus siue de pomis arborum Domini sunt et illi sanctifican-

27:31 And if a man will at all redeem ought of his tithes, tur 27:31 si quis autem uoluerit rediiere decimas suas addet

he shall add thereto the fifth part thereof. 27:32 And concerning the tithe of the herd, or of the flock, quintam partem earum 27:32 omnium decimarum boues et even of whatsoever passeth under the rod, the tenth shall be holy unto the LORD.

oues et caprae quae sub pastoris uirga transeunt quicquid

27:33 He shall not search whether it be good or bad, nei- decimum uenerit sanctificabitur Domino 27:33 non eligetur

ther shall he change it: and if he change it at all, then both it and the change thereof shall be holy; it shall not be re-

deemed. nec bonum nec malum nec altero commutabitur si quis mu-

Liber Leuiticus

tauerit et quod mutatum est et pro quo mutatum est sanctificabitur Domino et non rediietur 27:34 haec sunt praecepta 27:34 These are the commandments, which the LORD

commanded Moses for the children of Israel in mount Sinai.

quae mandauit Dominus Mosi ad filios Israhel in monte Sinai

Biblia Sacra Vulgata

CCXVI

\underline{Liber Numeri}

1:1 locutusque est Dominus ad Mosen in deserto Sinai 1:1 And the LORD spake unto Moses in the wilderness

of Sinai, in the tabernacle of the congregation, on the first day of the second month, in the second year after they

in tabernaculo foederis prima die mensis secundi anno al- were come out of the land of Egypt, saying, tero egressionis eorum ex Aegypto dicens 1:2 tollite sum- 1:2 Take ye the sum of all the congregation of the children

of Israel, after their families, by the house of their fathers, with the number of their names, every male by their polls;

mam uniuersae congregationis filiorum Israhel per cognationes et domos suas et nomina singulorum quicquid sexus est masculini 1:3 a uicesimo anno et supra omnium uiro- 1:3 From twenty years old and upward, all that are able

to go forth to war in Israel: thou and Aaron shall number them by their armies.

rum fortium ex Israhel et numerabitis eos per turmas suas tu et Aaron 1:4 eruntque uobiscum principes tribuum ac do- 1:4 And with you there shall be a man of every tribe; every

one head of the house of his fathers.

morum in cognationibus suis 1:5 quorum ista sunt nomina 1:5 And these are the names of the men that shall stand

with you: of the tribe of Reuben; Elizur the son of Shedeur.

de Ruben Elisur filius Sedeur 1:6 de Symeon Salamihel fi-

1:6 Of Simeon; Shelumiel the son of Zurishaddai.

lius Surisaddai 1:7 de Iuda Naasson filius Aminadab 1:8 1:7 Of Judah; Nahshon the son of Amminadab.

1:8 Of Issachar; Nethaneel the son of Zuar.

de Isachar Nathanahel filius Suar 1:9 de Zabulon Heliab _{1:9 Of Zebulun; Eliab the son of Helon.} filius Helon 1:10 filiorum autem Ioseph de Ephraim Heli- 1:10 Of the children of Joseph: of Ephraim; Elishama

the son of Ammihud: of Manasseh; Gamaliel the son of Pedahzur.

sama filius Ammiud de Manasse Gamalihel filius Phadassur 1:11 de Beniamin Abidan filius Gedeonis 1:12 de Dan 1:11 Of Benjamin; Abidan the son of Gideoni.

1:12 Of Dan; Ahiezer the son of Ammishaddai.

Ahiezer filius Amisaddai 1:13 de Aser Phegihel filius Och- _{1:13 Of Asher; Pagiel the son of Ocran.} ran 1:14 de Gad Heliasaph filius Duhel 1:15 de Nepthali 1:14 Of Gad; Eliasaph the son of Deuel.

1:15 Of Naphtali; Ahira the son of Enan.

Ahira filius Henan 1:16 hii nobilissimi principes multitudi- _{1:16 These were the renowned of the congregation,}

princes of the tribes of their fathers, heads of thousands in Israel.

nis per tribus et cognationes suas et capita exercitus Israhel 1:17 quos tulerunt Moses et Aaron cum omni uulgi multitu- 1:17 And Moses and Aaron took these men which are ex-

pressed by their names:

dine 1:18 et congregauerunt primo die mensis secundi re- 1:18 And they assembled all the congregation together on

the first day of the second month, and they declared their pedigrees after their families, by the house of their fathers, according to the number of the names, from twenty years old and upward, by their polls.

censentes eos per cognationes et domos ac familias et capita et nomina singulorum a uicesimo anno et supra 1:19 sicut 1:19 As the LORD commanded Moses, so he numbered

them in the wilderness of Sinai.

praeceperat Dominus Mosi numeratique sunt in deserto Sinai 1:20 de Ruben primogenito Israhelis per generationes 1:20 And the children of Reuben, Israel’s eldest son, by

their generations, after their families, by the house of their fathers, according to the number of the names, by their polls, every male from twenty years old and upward, all that were able to go forth to war;

et familias ac domos suas et nomina capitum singulorum omne quod sexus est masculini a uicesimo anno et supra procedentium ad bellum 1:21 quadraginta sex milia quin- 1:21 Those that were numbered of them, even of the tribe

of Reuben, were forty and six thousand and five hundred.

genti 1:22 de filiis Symeon per generationes et familias ac 1:22 Of the children of Simeon, by their generations, af-

ter their families, by the house of their fathers, those that were numbered of them, according to the number of the names, by their polls, every male from twenty years old and upward, all that were able to go forth to war;

domos cognationum suarum recensiti sunt per nomina et capita singulorum omne quod sexus est masculini a uicesimo anno et supra procedentium ad bellum 1:23 quinquaginta 1:23 Those that were numbered of them, even of the tribe

of Simeon, were fifty and nine thousand and three hundred.

nouem milia trecenti 1:24 de filiis Gad per generationes et

1:24 Of the children of Gad, by their generations, after their families, by the house of their fathers, according to the number of the names, from twenty years old and upward, all that were able to go forth to war;

Biblia Sacra Vulgata

familias ac domos cognationum suarum recensiti sunt per nomina singulorum a uiginti annis et supra omnes qui ad

1:25 Those that were numbered of them, even of the tribe bella procederent 1:25 quadraginta quinque milia sescenti

of Gad, were forty and five thousand six hundred and fifty. 1:26 Of the children of Judah, by their generations, after quinquaginta 1:26 de filiis Iuda per generationes et famitheir families, by the house of their fathers, according to the number of the names, from twenty years old and up-

ward, all that were able to go forth to war; lias ac domos cognationum suarum per nomina singulorum

a uicesimo anno et supra omnes qui poterant ad bella pro-

1:27 Those that were numbered of them, even of the tribe cedere 1:27 recensiti sunt septuaginta quattuor milia ses-

of Judah, were threescore and fourteen thousand and six hundred.

centi 1:28 de filiis Isachar per generationes et familias ac

1:28 Of the children of Issachar, by their generations, after their families, by the house of their fathers, according to the number of the names, from twenty years old and upward, all that were able to go forth to war;

domos cognationum suarum per nomina singulorum a uice-

1:29 Those that were numbered of them, even of the tribe simo anno et supra omnes qui ad bella procederent 1:29 re-

of Issachar, were fifty and four thousand and four hundred.

1:30 Of the children of Zebulun, by their generations, af- censiti sunt quinquaginta quattuor milia quadringenti 1:30

ter their families, by the house of their fathers, according to the number of the names, from twenty years old and

upward, all that were able to go forth to war; de filiis Zabulon per generationes et familias ac domos cognationum suarum recensiti sunt per nomina singulorum a uicesimo anno et supra omnes qui poterant ad bella procedere

1:31 Those that were numbered of them, even of the tribe 1:31 quinquaginta septem milia quadringenti 1:32 de filiis

of Zebulun, were fifty and seven thousand and four hundred. 1:32 Of the children of Joseph, namely, of the children of Ephraim, by their generations, after their families, by the house of their fathers, according to the number of the names, from twenty years old and upward, all that were able to go forth to war;

Ioseph filiorum Ephraim per generationes et familias ac domos cognationum suarum recensiti sunt per nomina singulorum a uicesimo anno et supra omnes qui poterant ad bella

1:33 Those that were numbered of them, even of the tribe procedere 1:33 quadraginta milia quingenti 1:34 porro fi-

of Ephraim, were forty thousand and five hundred. 1:34 Of the children of Manasseh, by their generations, after their families, by the house of their fathers, according to the number of the names, from twenty years old and upward, all that were able to go forth to war;

liorum Manasse per generationes et familias ac domos cognationum suarum recensiti sunt per nomina singulorum a uiginti annis et supra omnes qui poterant ad bella procedere

1:35 Those that were numbered of them, even of the tribe 1:35 triginta duo milia ducenti 1:36 de filiis Beniamin per

of Manasseh, were thirty and two thousand and two hundred. 1:36 Of the children of Benjamin, by their generations, after their families, by the house of their fathers, according to the number of the names, from twenty years old and upward, all that were able to go forth to war;

generationes et familias ac domos cognationum suarum recensiti sunt nominibus singulorum a uicesimo anno et supra

1:37 Those that were numbered of them, even of the tribe omnes qui poterant ad bella procedere 1:37 triginta quin-

of Benjamin, were thirty and five thousand and four hundred.

que milia quadringenti 1:38 de filiis Dan per generationes

1:38 Of the children of Dan, by their generations, after their families, by the house of their fathers, according to the number of the names, from twenty years old and upward, all that were able to go forth to war;

et familias ac domos cognationum suarum recensiti sunt nominibus singulorum a uicesimo anno et supra omnes qui po-

1:39 Those that were numbered of them, even of the tribe terant ad bella procedere 1:39 sexaginta duo milia septin-

of Dan, were threescore and two thousand and seven hundred.

genti 1:40 de filiis Aser per generationes et familias ac do-

1:40 Of the children of Asher, by their generations, after their families, by the house of their fathers, according to the number of the names, from twenty years old and upward, all that were able to go forth to war;

mos cognationum suarum recensiti sunt per nomina singulorum a uicesimo anno et supra omnes qui poterant ad bella procedere 1:41 quadraginta milia et mille quingenti 1:42 1:41 Those that were numbered of them, even of the tribe

of Asher, were forty and one thousand and five hundred. 1:42 Of the children of Naphtali, throughout their generations, after their families, by the house of their fathers, according to the number of the names, from twenty years old and upward, all that were able to go forth to war;

de filiis Nepthali per generationes et familias ac domos cognationum suarum recensiti sunt nominibus singulorum a uicesimo anno et supra omnes qui poterant ad bella procedere 1:43 quinquaginta tria milia quadringenti 1:44 hii sunt quos 1:43 Those that were numbered of them, even of the tribe

of Naphtali, were fifty and three thousand and four hundred. 1:44 These are those that were numbered, which Moses and Aaron numbered, and the princes of Israel, being twelve men: each one was for the house of his fathers.

numerauerunt Moses et Aaron et duodecim principes Israhel singulos per domos cognationum suarum 1:45 fueruntque

1:45 So were all those that were numbered of the children of Israel, by the house of their fathers, from twenty years old and upward, all that were able to go forth to war in Israel;

omnes filiorum Israhel per domos et familias suas a uicesimo anno et supra qui poterant ad bella procedere 1:46 sescenta 1:46 Even all they that were numbered were six hundred

thousand and three thousand and five hundred and fifty.

tria milia uirorum quingenti quinquaginta 1:47 Leuitae au- 1:47 But the Levites after the tribe of their fathers were

not numbered among them.

tem in tribu familiarum suarum non sunt numerati cum eis 1:48 locutusque est Dominus ad Mosen dicens 1:49 tribum 1:48 For the LORD had spoken unto Moses, saying,

1:49 Only thou shalt not number the tribe of Levi, neither take the sum of them among the children of Israel:

Leui noli numerare neque ponas summam eorum cum filiis Israhel 1:50 sed constitue eos super tabernaculum testimo- 1:50 But thou shalt appoint the Levites over the tabernacle

of testimony, and over all the vessels thereof, and over all things that belong to it: they shall bear the tabernacle, and all the vessels thereof; and they shall minister unto it, and shall encamp round about the tabernacle.

nii cuncta uasa eius et quicquid ad caerimonias pertinet ipsi portabunt tabernaculum et omnia utensilia eius et erunt in

ministerio ac per gyrum tabernaculi metabuntur 1:51 cum 1:51 And when the tabernacle setteth forward, the Levites shall take it down: and when the tabernacle is to be pitched, the Levites shall set it up: and the stranger that

proficiscendum fuerit deponent Leuitae tabernaculum cum cometh nigh shall be put to death. castra metanda erigent quisquis externorum accesserit occi-

detur 1:52 metabuntur autem castra filii Israhel unusquis- 1:52 And the children of Israel shall pitch their tents, every man by his own camp, and every man by his own standard, throughout their hosts.

que per turmas et cuneos atque exercitum suum 1:53 porro

1:53 But the Levites shall pitch round about the tabernacle of testimony, that there be no wrath upon the congregation of the children of Israel: and the Levites shall keep the charge of the tabernacle of testimony.

Leuitae per gyrum tabernaculi figent tentoria ne fiat indignatio super multitudinem filiorum Israhel et excubabunt in custodiis tabernaculi testimonii 1:54 fecerunt ergo filii Is- 1:54 And the children of Israel did according to all that the

LORD commanded Moses, so did they.

rahel iuxta omnia quae praeceperat Dominus Mosi

2:1 locutusque est Dominus ad Mosen et Aaron dicens 2:1 And the LORD spake unto Moses and unto Aaron,

saying,

2:2 singuli per turmas signa atque uexilla et domos cogna- 2:2 Every man of the children of Israel shall pitch by

his own standard, with the ensign of their father’s house: far off about the tabernacle of the congregation shall they

tionum suarum castrametabuntur filiorum Israhel per gyrum pitch. tabernaculi foederis 2:3 ad orientem Iudas figet tentoria per 2:3 And on the east side toward the rising of the sun shall

they of the standard of the camp of Judah pitch throughout their armies: and Nahshon the son of Amminadab shall be

turmas exercitus sui eritque princeps filiorum eius Naasson captain of the children of Judah.

filius Aminadab 2:4 et omnis de stirpe eius summa pug- 2:4 And his host, and those that were numbered of them,

were threescore and fourteen thousand and six hundred.

nantium septuaginta quattuor milia sescentorum 2:5 iuxta 2:5 And those that do pitch next unto him shall be the tribe

of Issachar: and Nethaneel the son of Zuar shall be captain of the children of Issachar.

eum castrametati sunt de tribu Isachar quorum princeps fuit

Biblia Sacra Vulgata

2:6 And his host, and those that were numbered thereof, Nathanahel filius Suar 2:6 et omnis numerus pugnatorum were fifty and four thousand and four hundred.

2:7 Then the tribe of Zebulun: and Eliab the son of Helon eius quinquaginta quattuor milia quadringenti 2:7 in tribu

shall be captain of the children of Zebulun. 2:8 And his host, and those that were numbered thereof, Zabulon princeps fuit Heliab filius Helon 2:8 omnis de were fifty and seven thousand and four hundred.

stirpe eius exercitus pugnatorum quinquaginta septem milia

2:9 All that were numbered in the camp of Judah were an quadringenti 2:9 uniuersi qui in castris Iudae adnumerati

hundred thousand and fourscore thousand and six thousand and four hundred, throughout their armies. These

shall first set forth. sunt fuerunt centum octoginta sex milia quadringenti et per 2:10 On the south side shall be the standard of the camp of turmas suas primi egredientur 2:10 in castris filiorum Ru-

Reuben according to their armies: and the captain of the children of Reuben shall be Elizur the son of Shedeur.

ben ad meridianam plagam erit princeps Elisur filius Sedeur

2:11 And his host, and those that were numbered thereof, 2:11 et cunctus exercitus pugnatorum eius qui numerati sunt

were forty and six thousand and five hundred.

2:12 And those which pitch by him shall be the tribe of quadraginta sex milia quingenti 2:12 iuxta eum castrame-

Simeon: and the captain of the children of Simeon shall be Shelumiel the son of Zurishaddai.

tati sunt de tribu Symeon quorum princeps fuit Salamihel

2:13 And his host, and those that were numbered of them, filius Surisaddai 2:13 et cunctus exercitus pugnatorum eius

were fifty and nine thousand and three hundred.

2:14 Then the tribe of Gad: and the captain of the sons of qui numerati sunt quinquaginta nouem milia trecenti 2:14

Gad shall be Eliasaph the son of Reuel.

2:15 And his host, and those that were numbered of them, in tribu Gad princeps fuit Heliasaph filius Duhel 2:15 et

were forty and five thousand and six hundred and fifty.

cunctus exercitus pugnatorum eius qui numerati sunt qua-

2:16 All that were numbered in the camp of Reuben were draginta quinque milia sescenti quinquaginta 2:16 omnes

an hundred thousand and fifty and one thousand and four hundred and fifty, throughout their armies. And they shall

set forth in the second rank. qui recensiti sunt in castris Ruben centum quinquaginta mi-

lia et mille quadringenti quinquaginta per turmas suas in se-

2:17 Then the tabernacle of the congregation shall set for- cundo loco proficiscentur 2:17 leuabitur autem tabernacuward with the camp of the Levites in the midst of the camp: as they encamp, so shall they set forward, every

man in his place by their standards. lum testimonii per officia Leuitarum et turmas eorum quo-

modo erigetur ita et deponetur singuli per loca et ordines

2:18 On the west side shall be the standard of the camp of suos proficiscentur 2:18 ad occidentalem plagam erunt cas-

Ephraim according to their armies: and the captain of the sons of Ephraim shall be Elishama the son of Ammihud.

tra filiorum Ephraim quorum princeps fuit Helisama filius

2:19 And his host, and those that were numbered of them, Ammiud 2:19 cunctus exercitus pugnatorum eius qui nuwere forty thousand and five hundred.

2:20 And by him shall be the tribe of Manasseh: and the merati sunt quadraginta milia quingenti 2:20 et cum eis tri-

captain of the children of Manasseh shall be Gamaliel the son of Pedahzur.

bus filiorum Manasse quorum princeps fuit Gamalihel filius

2:21 And his host, and those that were numbered of them, Phadassur 2:21 cunctus exercitus pugnatorum eius qui nu-

were thirty and two thousand and two hundred.

2:22 Then the tribe of Benjamin: and the captain of the merati sunt triginta duo milia ducenti 2:22 in tribu filio-

sons of Benjamin shall be Abidan the son of Gideoni.

2:23 And his host, and those that were numbered of them, rum Beniamin princeps fuit Abidan filius Gedeonis 2:23 et

were thirty and five thousand and four hundred.

cunctus exercitus pugnatorum eius qui recensiti sunt triginta

2:24 All that were numbered of the camp of Ephraim were quinque milia quadringenti 2:24 omnes qui numerati sunt

an hundred thousand and eight thousand and an hundred, throughout their armies. And they shall go forward in the

third rank. in castris Ephraim centum octo milia centum per turmas suas 2:25 The standard of the camp of Dan shall be on the north tertii proficiscentur 2:25 ad aquilonis partem castrametati

side by their armies: and the captain of the children of Dan shall be Ahiezer the son of Ammishaddai.

sunt filii Dan quorum princeps fuit Ahiezer filius Amisaddai 2:26 cunctus exercitus pugnatorum eius qui numerati 2:26 And his host, and those that were numbered of them,

were threescore and two thousand and seven hundred.

sunt sexaginta duo milia septingenti 2:27 iuxta eum fixere 2:27 And those that encamp by him shall be the tribe of

Asher: and the captain of the children of Asher shall be Pagiel the son of Ocran.

tentoria de tribu Aser quorum princeps fuit Phegihel filius Ochran 2:28 cunctus exercitus pugnatorum eius qui nume- 2:28 And his host, and those that were numbered of them,

were forty and one thousand and five hundred.

rati sunt quadraginta milia et mille quingenti 2:29 de tribu 2:29 Then the tribe of Naphtali: and the captain of the

children of Naphtali shall be Ahira the son of Enan.

filiorum Nepthalim princeps fuit Ahira filius Henan 2:30 2:30 And his host, and those that were numbered of them,

were fifty and three thousand and four hundred.

cunctus exercitus pugnatorum eius quinquaginta tria milia quadringenti 2:31 omnes qui numerati sunt in castris Dan 2:31 All they that were numbered in the camp of Dan were

an hundred thousand and fifty and seven thousand and six hundred. They shall go hindmost with their standards.

fuerunt centum quinquaginta septem milia sescenti et nouissimi proficiscentur 2:32 hic numerus filiorum Israhel per 2:32 These are those which were numbered of the children

of Israel by the house of their fathers: all those that were numbered of the camps throughout their hosts were six hundred thousand and three thousand and five hundred and fifty.

domos cognationum suarum et turmas diuisi exercitus sescenta tria milia quingenti quinquaginta 2:33 Leuitae autem 2:33 But the Levites were not numbered among the chil-

dren of Israel; as the LORD commanded Moses.

non sunt numerati inter filios Israhel sic enim praecepit Dominus Mosi 2:34 feceruntque filii Israhel iuxta omnia quae 2:34 And the children of Israel did according to all that

the LORD commanded Moses: so they pitched by their standards, and so they set forward, every one after their

mandauerat Dominus castrametati sunt per turmas suas et families, according to the house of their fathers. profecti per familias ac domos patrum suorum

3:1 haec sunt generationes Aaron et Mosi in die qua lo- 3:1 These also are the generations of Aaron and Moses in

the day that the LORD spake with Moses in mount Sinai.

cutus est Dominus ad Mosen in monte Sinai 3:2 et haec 3:2 And these are the names of the sons of Aaron; Nadab

the firstborn, and Abihu, Eleazar, and Ithamar.

nomina filiorum Aaron primogenitus eius Nadab dein Abiu et Eleazar et Ithamar 3:3 haec nomina filiorum Aaron sa- 3:3 These are the names of the sons of Aaron, the priests

which were anointed, whom he consecrated to minister in the priest’s office.

cerdotum qui uncti sunt et quorum repletae et consecratae manus ut sacerdotio fungerentur 3:4 mortui sunt Nadab et 3:4 And Nadab and Abihu died before the LORD, when

they offered strange fire before the LORD, in the wilderness of Sinai, and they had no children: and Eleazar and Ithamar ministered in the priest’s office in the sight of Aaron their father.

Abiu cum offerrent ignem alienum in conspectu Domini in deserto Sinai absque liberis functique sunt sacerdotio Eleazar et Ithamar coram Aaron patre suo 3:5 locutus est Do- 3:5 And the LORD spake unto Moses, saying, minus ad Mosen dicens 3:6 adplica tribum Leui et fac stare 3:6 Bring the tribe of Levi near, and present them before

Aaron the priest, that they may minister unto him.

in conspectu Aaron sacerdotis ut ministrent ei et excubent 3:7 et obseruent quicquid ad cultum pertinet multitudinis co- 3:7 And they shall keep his charge, and the charge of the

whole congregation before the tabernacle of the congregation, to do the service of the tabernacle.

ram tabernaculo testimonii 3:8 et custodiant uasa taberna-

3:8 And they shall keep all the instruments of the tabernacle of the congregation, and the charge of the children of Israel, to do the service of the tabernacle.

culi seruientes in ministerio eius 3:9 dabisque dono Leuitas

3:9 And thou shalt give the Levites unto Aaron and to his sons: they are wholly given unto him out of the children

3:10 Aaron et filiis eius quibus traditi sunt a filiis Israhel Aa- _{of Israel.}

3:10 And thou shalt appoint Aaron and his sons, and they shall wait on their priest’s office: and the stranger that cometh nigh shall be put to death.

ron autem et filios eius constitues super cultum sacerdotii 3:11 And the LORD spake unto Moses, saying, externus qui ad ministrandum accesserit morietur 3:11 lo-

3:12 And I, behold, I have taken the Levites from among cutusque est Dominus ad Mosen dicens 3:12 ego tuli Lethe children of Israel instead of all the firstborn that openeth the matrix among the children of Israel: therefore

the Levites shall be mine; uitas a filiis Israhel pro omni primogenito qui aperit uul-

3:13 Because all the firstborn are mine; for on the day that uam in filiis Israhel eruntque Leuitae mei 3:13 meum est I smote all the firstborn in the land of Egypt I hallowed unto me all the firstborn in Israel, both man and beast:

mine shall they be: I am the LORD. enim omne primogenitum ex quo percussi primogenitos in terra Aegypti sanctificaui mihi quicquid primum nascitur in Israhel ab homine usque ad pecus mei sunt ego Dominus

3:14 And the LORD spake unto Moses in the wilderness 3:14 locutus est Dominus ad Mosen in deserto Sinai dicens

of Sinai, saying,

3:15 Number the children of Levi after the house of their 3:15 numera filios Leui per domos patrum suorum et famifathers, by their families: every male from a month old and upward shalt thou number them.

lias omnem masculum ab uno mense et supra 3:16 numera-

3:16 And Moses numbered them according to the word of the LORD, as he was commanded.

3:17 And these were the sons of Levi by their names; Ger- uit Moses ut praeceperat Dominus 3:17 et inuenti sunt filii

shon, and Kohath, and Merari.

3:18 And these are the names of the sons of Gershon by Leui per nomina sua Gerson et Caath et Merari 3:18 filii

their families; Libni, and Shimei.

3:19 And the sons of Kohath by their families; Amram, Gerson Lebni et Semei 3:19 filii Caath Amram et Iessaar

and Izehar, Hebron, and Uzziel.

3:20 And the sons of Merari by their families; Mahli, and Hebron et Ozihel 3:20 filii Merari Mooli et Musi 3:21 de Mushi. These are the families of the Levites according to the house of their fathers.

3:21 Of Gershon was the family of the Libnites, and the family of the Shimites: these are the families of the Gershonites.

Gerson fuere familiae duae lebnitica et semeitica 3:22 qua-

3:22 Those that were numbered of them, according to the number of all the males, from a month old and upward, even those that were numbered of them were seven thousand and five hundred.

rum numeratus est populus sexus masculini ab uno mense et supra septem milia quingentorum 3:23 hii post tabernacu-

3:23 The families of the Gershonites shall pitch behind the

tabernacle westward. lum metabuntur ad occidentem 3:24 sub principe Eliasaph

3:24 And the chief of the house of the father of the Gershonites shall be Eliasaph the son of Lael.

filio Lahel 3:25 et habebunt excubias in tabernaculo foe-

3:25 And the charge of the sons of Gershon in the tabernacle of the congregation shall be the tabernacle, and the tent, the covering thereof, and the hanging for the door of the tabernacle of the congregation,

deris 3:26 ipsum tabernaculum et operimentum eius tento-

3:26 And the hangings of the court, and the curtain for the door of the court, which is by the tabernacle, and by the altar round about, and the cords of it for all the service thereof.

rium quod trahitur ante fores tecti foederis et cortinas atrii tentorium quoque quod adpenditur in introitu atrii tabernaculi et quicquid ad ritum altaris pertinet funes tabernaculi et

3:27 And of Kohath was the family of the Amramites, and omnia utensilia eius 3:27 cognatio Caath habebit populos

the family of the Izeharites, and the family of the Hebronites, and the family of the Uzzielites: these are the

families of the Kohathites. Amramitas et Iessaaritas et Hebronitas et Ozihelitas hae sunt 3:28 In the number of all the males, from a month old and familiae Caathitarum recensitae per nomina sua 3:28 om-

upward, were eight thousand and six hundred, keeping the charge of the sanctuary.

nes generis masculini ab uno mense et supra octo milia se-

3:29 The families of the sons of Kohath shall pitch on the scenti habebunt excubias sanctuarii 3:29 et castrametabun-

side of the tabernacle southward.

3:30 And the chief of the house of the father of the families tur ad meridianam plagam 3:30 princepsque eorum erit Eliof the Kohathites shall be Elizaphan the son of Uzziel.

3:31 And their charge shall be the ark, and the table, and saphan filius Ozihel 3:31 et custodient arcam mensamque the candlestick, and the altars, and the vessels of the sanctuary wherewith they minister, and the hanging, and all the

service thereof. et candelabrum altaria et uasa sanctuarii in quibus ministra-

3:32 And Eleazar the son of Aaron the priest shall be chief tur et uelum cunctamque huiuscemodi supellectilem 3:32 over the chief of the Levites, and have the oversight of them that keep the charge of the sanctuary.

princeps autem principum Leuitarum Eleazar filius Aaron sacerdotis erit super excubitores custodiae sanctuarii 3:33 3:33 Of Merari was the family of the Mahlites, and the

family of the Mushites: these are the families of Merari.

at uero de Merari erunt populi Moolitae et Musitae recensiti per nomina sua 3:34 omnes generis masculini ab uno 3:34 And those that were numbered of them, according to

the number of all the males, from a month old and upward, were six thousand and two hundred.

mense et supra sex milia ducenti 3:35 princeps eorum Suri-

3:35 And the chief of the house of the father of the families of Merari was Zuriel the son of Abihail: these shall pitch on the side of the tabernacle northward.

hel filius Abiahihel in plaga septentrionali castrametabuntur 3:36 erunt sub custodia eorum tabulae tabernaculi et uectes et 3:36 And under the custody and charge of the sons of

Merari shall be the boards of the tabernacle, and the bars thereof, and the pillars thereof, and the sockets thereof,

columnae ac bases earum et omnia quae ad cultum huiusce- and all the vessels thereof, and all that serveth thereto,

modi pertinent 3:37 columnaeque atrii per circuitum cum 3:37 And the pillars of the court round about, and their

sockets, and their pins, and their cords.

basibus suis et paxilli cum funibus 3:38 castrametabuntur 3:38 But those that encamp before the tabernacle toward

the east, even before the tabernacle of the congregation eastward, shall be Moses, and Aaron and his sons, keeping the charge of the sanctuary for the charge of the children of Israel; and the stranger that cometh nigh shall be put to death.

ante tabernaculum foederis id est ad orientalem plagam Moses et Aaron cum filiis suis habentes custodiam sanctuarii in medio filiorum Israhel quisquis alienus accesserit morietur 3:39 omnes Leuitae quos numerauerunt Moses et Aaron 3:39 All that were numbered of the Levites, which Moses

and Aaron numbered at the commandment of the LORD, throughout their families, all the males from a month old

iuxta praeceptum Domini per familias suas in genere mascu- and upward, were twenty and two thousand. lino a mense uno et supra fuerunt uiginti duo milia 3:40 et 3:40 And the LORD said unto Moses, Number all the first-

born of the males of the children of Israel from a month old and upward, and take the number of their names.

ait Dominus ad Mosen numera primogenitos sexus masculini de filiis Israhel a mense uno et supra et habebis summam eorum 3:41 tollesque Leuitas mihi pro omni primogenito 3:41 And thou shalt take the Levites for me (I am the

LORD) instead of all the firstborn among the children of Israel; and the cattle of the Levites instead of all the

filiorum Israhel ego sum Dominus et pecora eorum pro uni- firstlings among the cattle of the children of Israel. uersis primogenitis pecoris filiorum Israhel 3:42 recensuit 3:42 And Moses numbered, as the LORD commanded

him, all the firstborn among the children of Israel.

Moses sicut praeceperat Dominus primogenitos filiorum Israhel 3:43 et fuerunt masculi per nomina sua a mense uno 3:43 And all the firstborn males by the number of names,

from a month old and upward, of those that were numbered of them, were twenty and two thousand two hundred

et supra uiginti duo milia ducenti septuaginta tres 3:44 lo- and threescore and thirteen.

3:44 And the LORD spake unto Moses, saying,

cutusque est Dominus ad Mosen 3:45 tolle Leuitas pro pri- 3:45 Take the Levites instead of all the firstborn among the children of Israel, and the cattle of the Levites instead of their cattle; and the Levites shall be mine: I am the LORD.

mogenitis filiorum Israhel et pecora Leuitarum pro pecori-

bus eorum eruntque Leuitae mei ego sum Dominus 3:46 in 3:46 And for those that are to be redeemed of the two hundred and threescore and thirteen of the firstborn of the children of Israel, which are more than the Levites;

pretio autem ducentorum septuaginta trium qui excedunt numerum Leuitarum de primogenitis filiorum Israhel 3:47 ac- 3:47 Thou shalt even take five shekels apiece by the poll,

after the shekel of the sanctuary shalt thou take them: (the shekel is twenty gerahs:)

cipies quinque siclos per singula capita ad mensuram sanctuarii siclus habet obolos uiginti 3:48 dabisque pecuniam 3:48 And thou shalt give the money, wherewith the odd

number of them is to be redeemed, unto Aaron and to his sons.

Aaron et filiis eius pretium eorum qui supra sunt 3:49 tulit

3:49 And Moses took the redemption money of them that were over and above them that were redeemed by the Levites:

igitur Moses pecuniam eorum qui fuerant amplius et quos redeeerant a Leuitis 3:50 pro primogenitis filiorum Israhel 3:50 Of the firstborn of the children of Israel took he the

money; a thousand three hundred and threescore and five shekels, after the shekel of the sanctuary:

Biblia Sacra Vulgata

mille trecentorum sexaginta quinque siclorum iuxta pondus

3:51 And Moses gave the money of them that were re- sanctuarii 3:51 et dedit eam Aaroni et filiis eius iuxta uer-

deemed unto Aaron and to his sons, according to the word of the LORD, as the LORD commanded Moses.

bum quod praeceperat sibi Dominus

4:1 And the LORD spake unto Moses and unto Aaron, 4:1 locutusque est Dominus ad Mosen et Aaron dicens

saying,

4:2 Take the sum of the sons of Kohath from among the 4:2 tolle summam filiorum Caath de medio Leuitarum per

sons of Levi, after their families, by the house of their fathers,

domos et familias suas 4:3 a tricesimo anno et supra us-

4:3 From thirty years old and upward even until fifty years old, all that enter into the host, to do the work in the tabernacle of the congregation.

que ad quinquagesimum annum omnium qui ingrediuntur ut

4:4 This shall be the service of the sons of Kohath in the stent et ministrent in tabernaculo foederis 4:4 hic est cultus

tabernacle of the congregation, about the most holy things:

filiorum Caath tabernaculum foederis et sanctum sanctorum

4:5 And when the camp setteth forward, Aaron shall come, 4:5 ingredientur Aaron et filii eius quando mouenda sunt cas-

and his sons, and they shall take down the covering vail, and cover the ark of testimony with it:

tra et deponent uelum quod pendet ante fores inuoluentque

4:6 And shall put thereon the covering of badgers’ skins, eo arcam testimonii 4:6 et operient rursum uelamine iant-

and shall spread over it a cloth wholly of blue, and shall put in the staves thereof.

hinarum pellium extendentque desuper pallium totum hya-

4:7 And upon the table of shewbread they shall spread a cinthinum et inducent uectes 4:7 mensam quoque propo-

cloth of blue, and put thereon the dishes, and the spoons, and the bowls, and covers to cover withal: and the contin-

ual bread shall be thereon: sitionis inuoluent hyacinthino pallio et ponent cum ea turi-

bula et mortariola cyatos et crateras ad liba fundenda panes

4:8 And they shall spread upon them a cloth of scarlet, and semper in ea erunt 4:8 extendentque desuper pallium coccover the same with a covering of badgers’ skins, and shall put in the staves thereof.

cineum quod rursum operient uelamento ianthinarum pel-

4:9 And they shall take a cloth of blue, and cover the can- lium et inducent uectes 4:9 sument et pallium hyacinthi-

dlestick of the light, and his lamps, and his tongs, and his snuffdishes, and all the oil vessels thereof, wherewith they

minister unto it: num quo operient candelabrum cum lucernis et forcipibus

suis et emunctoriis et cunctis uasis olei quae ad concinnan-

4:10 And they shall put it and all the vessels thereof within das lucernas necessaria sunt 4:10 et super omnia ponent a covering of badgers’ skins, and shall put it upon a bar.

4:11 And upon the golden altar they shall spread a cloth of operimentum ianthinarum pellium et inducent uectes 4:11

blue, and cover it with a covering of badgers’ skins, and shall put to the staves thereof:

nec non et altare aureum inuoluent hyacinthino uestimento et extendent desuper operimentum ianthinarum pellium indu-

4:12 And they shall take all the instruments of ministry, centque uectes 4:12 omnia uasa quibus ministratur in sanc-

wherewith they minister in the sanctuary, and put them in a cloth of blue, and cover them with a covering of badgers’

skins, and shall put them on a bar: tuario inuoluent hyacinthino pallio et extendent desuper ope-4:13 And they shall take away the ashes from the altar, and rimentum ianthinarum pellium inducentque uectes 4:13 sed

spread a purple cloth thereon:

et altare mundabunt cinere et inuoluent illud purpureo ues-

4:14 And they shall put upon it all the vessels thereof, timento 4:14 ponentque cum eo omnia uasa quibus in miwherewith they minister about it, even the censers, the fleshhooks, and the shovels, and the basons, all the vessels of the altar; and they shall spread upon it a covering of badgers’ skins, and put to the staves of it.

nisterio eius utuntur id est ignium receptacula fuscinulas ac tridentes uncinos et uatilla cuncta uasa altaris operient simul

4:15 And when Aaron and his sons have made an end of uelamine ianthinarum pellium et inducent uectes 4:15 cum-

covering the sanctuary, and all the vessels of the sanctuary, as the camp is to set forward; after that, the sons of Kohath shall come to bear it: but they shall not touch any holy thing, lest they die. These things are the burden of the sons of Kohath in the tabernacle of the congregation.

CCXXIV

que inuoluerint Aaron et filii eius sanctuarium et omnia uasa eius in commotione castrorum tunc intrabunt filii Caath ut portent inuoluta et non tangant uasa sanctuarii ne moriantur ista sunt onera filiorum Caath in tabernaculo foederis 4:16 4:16 And to the office of Eleazar the son of Aaron the

priest pertaineth the oil for the light, and the sweet incense, and the daily meat offering, and the anointing oil, and the oversight of all the tabernacle, and of all that therein is, in the sanctuary, and in the vessels thereof.

super quos erit Eleazar filius Aaron sacerdotis ad cuius pertinet curam oleum ad concinnandas lucernas et conpositionis incensum et sacrificium quod semper offertur et oleum unctionis et quicquid ad cultum tabernaculi pertinet omniumque uasorum quae in sanctuario sunt 4:17 locutusque est Domi- 4:17 And the LORD spake unto Moses and unto Aaron

saying,

nus ad Mosen et Aaron dicens 4:18 nolite perdere populum 4:18 Cut ye not off the tribe of the families of the Ko-

hathites from among the Levites:

Caath de medio Leuitarum 4:19 sed hoc facite eis ut ui- 4:19 But thus do unto them, that they may live, and not

die, when they approach unto the most holy things: Aaron and his sons shall go in, and appoint them every one to his

uant et non moriantur si tetigerint sancta sanctorum Aaron service and to his burden: et filii eius intrabunt ipsique disponent opera singulorum et diuident quid portare quis debeat 4:20 alii nulla curiositate 4:20 But they shall not go in to see when the holy things

are covered, lest they die.

uideant quae sunt in sanctuario priusquam inuoluantur alioquin morientur 4:21 locutus est Dominus ad Mosen dicens 4:21 And the LORD spake unto Moses, saying,

4:22 tolle summam etiam filiorum Gerson per domos ac fa- 4:22 Take also the sum of the sons of Gershon, throughout

the houses of their fathers, by their families;

milias et cognationes suas 4:23 a triginta annis et supra us- 4:23 From thirty years old and upward until fifty years old

shalt thou number them; all that enter in to perform the service, to do the work in the tabernacle of the congrega-

que ad annos quinquaginta numera omnes qui ingrediuntur tion. et ministrant in tabernaculo foederis 4:24 hoc est officium 4:24 This is the service of the families of the Gershonites,

to serve, and for burdens:

familiae Gersonitarum 4:25 ut portent cortinas tabernaculi 4:25 And they shall bear the curtains of the tabernacle, and

the tabernacle of the congregation, his covering, and the covering of the badgers’ skins that is above upon it, and the

et tectum foederis operimentum aliud et super omnia uela- hanging for the door of the tabernacle of the congregation,

men ianthinum tentoriumque quod pendet in introitu foederis tabernaculi 4:26 cortinas atrii et uelum in introitu quod 4:26 And the hangings of the court, and the hanging for

the door of the gate of the court, which is by the tabernacle and by the altar round about, and their cords, and all the instruments of their service, and all that is made for them: so shall they serve.

est ante tabernaculum omnia quae ad altare pertinent funiculos et uasa ministerii 4:27 iubente Aaron et filiis eius por- 4:27 At the appointment of Aaron and his sons shall be

all the service of the sons of the Gershonites, in all their burdens, and in all their service: and ye shall appoint unto

tabunt filii Gerson et scient singuli cui debeant oneri man- them in charge all their burdens. cipari 4:28 hic est cultus familiae Gersonitarum in taberna- 4:28 This is the service of the families of the sons of

Gershon in the tabernacle of the congregation: and their charge shall be under the hand of Ithamar the son of Aaron

culo foederis eruntque sub manu Ithamar filii Aaron sacer- the priest. dotis 4:29 filios quoque Merari per familias et domos pat- 4:29 As for the sons of Merari, thou shalt number them

after their families, by the house of their fathers;

rum suorum recensebis 4:30 a triginta annis et supra usque 4:30 From thirty years old and upward even unto fifty

years old shalt thou number them, every one that entereth into the service, to do the work of the tabernacle of the

ad annos quinquaginta omnes qui ingrediuntur ad officium congregation. ministerii sui et cultum foederis testimonii 4:31 haec sunt 4:31 And this is the charge of their burden, according to

all their service in the tabernacle of the congregation; the boards of the tabernacle, and the bars thereof, and the pillars thereof, and sockets thereof,

Biblia Sacra Vulgata

onera eorum portabunt tabulas tabernaculi et uectes eius co-

4:32 And the pillars of the court round about, and their lumnas et bases earum 4:32 columnas quoque atrii per cir-

sockets, and their pins, and their cords, with all their instruments, and with all their service: and by name ye shall

reckon the instruments of the charge of their burden. cuitum cum basibus et paxillis et funibus suis omnia uasa et 4:33 This is the service of the families of the sons of Mer- supellectilem ad numerum accipient sicque portabunt 4:33

ari, according to all their service, in the tabernacle of the congregation, under the hand of Ithamar the son of Aaron

the priest. hoc est officium familiae Meraritarum et ministerium in ta-

bernaculo foederis eruntque sub manu Ithamar filii Aaron

4:34 And Moses and Aaron and the chief of the congrega- sacerdotis 4:34 recensuerunt igitur Moses et Aaron et prin-

tion numbered the sons of the Kohathites after their families, and after the house of their fathers,

cipes synagogae filios Caath per cognationes et domos pat-

4:35 From thirty years old and upward even unto fifty rum suorum 4:35 a triginta annis et supra usque ad annum

years old, every one that entereth into the service, for the work in the tabernacle of the congregation:

quinquagesimum omnes qui ingrediuntur ad ministerium ta-

4:36 And those that were numbered of them by their fam- bernaculi foederis 4:36 et inuenti sunt duo milia septingenti

ilies were two thousand seven hundred and fifty.

4:37 These were they that were numbered of the families quinquaginta 4:37 hic est numerus populi Caath qui intrat

of the Kohathites, all that might do service in the tabernacle of the congregation, which Moses and Aaron did number according to the commandment of the LORD by the hand of Moses.

tabernaculum foederis hos numerauit Moses et Aaron iuxta

4:38 And those that were numbered of the sons of Ger- sermonem Domini per manum Mosi 4:38 numerati sunt et

shon, throughout their families, and by the house of their fathers,

filii Gerson per cognationes et domos patrum suorum 4:39

4:39 From thirty years old and upward even unto fifty years old, every one that entereth into the service, for the work in the tabernacle of the congregation,

a triginta annis et supra usque ad annum quinquagesimum omnes qui ingrediuntur ut ministrent in tabernaculo foederis

4:40 Even those that were numbered of them, throughout 4:40 et inuenti sunt duo milia sescenti triginta 4:41 hic est

their families, by the house of their fathers, were two thousand and six hundred and thirty. 4:41 These are they that were numbered of the families of the sons of Gershon, of all that might do service in the tabernacle of the congregation, whom Moses and Aaron did number according to the commandment of the LORD.

populus Gersonitarum quos numerauerunt Moses et Aaron iuxta uerbum Domini 4:42 numerati sunt et filii Merari per

4:42 And those that were numbered of the families of the sons of Merari, throughout their families, by the house of their fathers,

cognationes et domos patrum suorum 4:43 a triginta annis

4:43 From thirty years old and upward even unto fifty years old, every one that entereth into the service, for the work in the tabernacle of the congregation,

et supra usque ad annum quinquagesimum omnes qui ingrediuntur ad explendos ritus tabernaculi foederis 4:44 et in-

4:44 Even those that were numbered of them after their families, were three thousand and two hundred.

4:45 These be those that were numbered of the families uenti sunt tria milia ducenti 4:45 hic est numerus filiorum

of the sons of Merari, whom Moses and Aaron numbered according to the word of the LORD by the hand of Moses.

Merari quos recensuerunt Moses et Aaron iuxta imperium

4:46 All those that were numbered of the Levites, whom Domini per manum Mosi 4:46 omnes qui recensiti sunt de Moses and Aaron and the chief of Israel numbered, after their families, and after the house of their fathers,

Leuitis et quos fecit ad nomen Moses et Aaron et princi-

4:47 From thirty years old and upward even unto fifty pes Israhel per cognationes et domos patrum suorum 4:47

years old, every one that came to do the service of the ministry, and the service of the burden in the tabernacle of

the congregation. a triginta annis et supra usque ad annum quinquagesimum

ingredientes ad ministerium tabernaculi et onera portanda

4:48 Even those that were numbered of them, were eight 4:48 fuerunt simul octo milia quingenti octoginta 4:49 iuxta

thousand and five hundred and fourscore, 4:49 According to the commandment of the LORD they were numbered by the hand of Moses, every one according to his service, and according to his burden: thus were they numbered of him, as the LORD commanded Moses.

uerbum Domini recensuit eos Moses unumquemque iuxta officium et onera sua sicut praeceperat ei Dominus 5:1 locutusque est Dominus ad Mosen dicens 5:2 prae- 5:1 And the LORD spake unto Moses, saying,

5:2 Command the children of Israel, that they put out of the camp every leper, and every one that hath an issue, and whosoever is defiled by the dead:

cipe filiis Israhel ut eiciant de castris omnem leprosum et qui semine fluit pollutusque est super mortuo 5:3 tam ma- 5:3 Both male and female shall ye put out, without the

camp shall ye put them; that they defile not their camps, in the midst whereof I dwell.

sculum quam feminam eicite de castris ne contaminent ea cum habitauerim uobiscum 5:4 feceruntque ita filii Israhel 5:4 And the children of Israel did so, and put them out

without the camp: as the LORD spake unto Moses, so did the children of Israel.

et eiecerunt eos extra castra sicut locutus erat Dominus Mosi 5:5 locutus est Dominus ad Mosen dicens 5:6 loquere ad fi- 5:5 And the LORD spake unto Moses, saying,

5:6 Speak unto the children of Israel, When a man or woman shall commit any sin that men commit, to do a trespass against the LORD, and that person be guilty;

lios Israhel uir siue mulier cum fecerint ex omnibus peccatis quae solent hominibus accidere et per neglegentiam transgressi fuerint mandatum Domini atque deliquerint 5:7 con- 5:7 Then they shall confess their sin which they have done:

and he shall recompense his trespass with the principal thereof, and add unto it the fifth part thereof, and give it

fitebuntur peccatum suum et reddent ipsum caput quintam- unto him against whom he hath trespassed. que partem desuper ei in quem peccauerint 5:8 sin autem 5:8 But if the man have no kinsman to recompense the

trespass unto, let the trespass be recompensed unto the LORD, even to the priest; beside the ram of the atone-

non fuerit qui recipiat dabunt Domino et erit sacerdotis ex- ment, whereby an atonement shall be made for him. cepto ariete qui offertur pro expiatione ut sit placabilis hostia 5:9 omnes quoque primitiae quas offerunt filii Israhel ad sa- 5:9 And every offering of all the holy things of the children

of Israel, which they bring unto the priest, shall be his.

cerdotem pertinent 5:10 et quicquid in sanctuarium offer- 5:10 And every man’s hallowed things shall be his: what-

soever any man giveth the priest, it shall be his.

tur a singulis et traditur manibus sacerdotis ipsius erit 5:11 5:11 And the LORD spake unto Moses, saying,

locutus est Dominus ad Mosen dicens 5:12 loquere ad fi- 5:12 Speak unto the children of Israel, and say unto them, If any man’s wife go aside, and commit a trespass against him,

lios Israhel et dices ad eos uir cuius uxor errauerit maritumque contemnens 5:13 dormierit cum altero uiro et hoc ma- 5:13 And a man lie with her carnally, and it be hid from the

eyes of her husband, and be kept close, and she be defiled, and there be no witness against her, neither she be taken

ritus deprehendere non quiuerit sed latet adulterium et testi- with the manner; bus argui non potest quia non est inuenta in stupro 5:14 si 5:14 And the spirit of jealousy come upon him, and he be

jealous of his wife, and she be defiled: or if the spirit of jealousy come upon him, and he be jealous of his wife,

spiritus zelotypiae concitauerit uirum contra uxorem suam and she be not defiled: quae uel polluta est uel falsa suspicione appetitur 5:15 ad- 5:15 Then shall the man bring his wife unto the priest,

and he shall bring her offering for her, the tenth part of an ephah of barley meal; he shall pour no oil upon it, nor put frankincense thereon; for it is an offering of jealousy, an offering of memorial, bringing iniquity to remembrance.

ducet eam ad sacerdotem et offeret oblationem pro illa decimam partem sati farinae hordiaciae non fundet super eam oleum nec inponet tus quia sacrificium zelotypiae est et oblatio inuestigans adulterium 5:16 offeret igitur eam sacerdos 5:16 And the priest shall bring her near, and set her before

the LORD:

et statuet coram Domino 5:17 adsumetque aquam sanctam 5:17 And the priest shall take holy water in an earthen

vessel; and of the dust that is in the floor of the tabernacle the priest shall take, and put it into the water:

in uase fictili et pauxillum terrae de pauimento tabernaculi mittet in eam 5:18 cumque steterit mulier in conspectu Do- 5:18 And the priest shall set the woman before the LORD,

and uncover the woman’s head, and put the offering of memorial in her hands, which is the jealousy offering: and the priest shall have in his hand the bitter water that causeth the curse:

mini discoperiet caput eius et ponet super manus illius sacrificium recordationis et oblationem zelotypiae ipse autem

tenebit aquas amarissimas in quibus cum execratione male-

5:19 And the priest shall charge her by an oath, and say dicta congessit 5:19 adiurabitque eam et dicet si non dor-

unto the woman, If no man have lain with thee, and if thou hast not gone aside to uncleanness with another instead of thy husband, be thou free from this bitter water that causeth the curse:

miuit uir alienus tecum et si non polluta es deserto mariti toro non te nocebunt aquae istae amarissimae in quas ma-

5:20 But if thou hast gone aside to another instead of thy ledicta congessi 5:20 sin autem declinasti a uiro tuo atque

husband, and if thou be defiled, and some man have lain with thee beside thine husband:

polluta es et concubuisti cum altero 5:21 his maledictioni-

5:21 Then the priest shall charge the woman with an oath of cursing, and the priest shall say unto the woman, The LORD make thee a curse and an oath among thy people, when the LORD doth make thy thigh to rot, and thy belly to swell;

bus subiacebis det te Dominus in maledictionem exemplumque cunctorum in populo suo putrescere faciat femur tuum

5:22 And this water that causeth the curse shall go into thy et tumens uterus disrumpatur 5:22 ingrediantur aquae mabowels, to make thy belly to swell, and thy thigh to rot: And the woman shall say, Amen, amen.

ledictae in uentrem tuum et utero tumescente putrescat fe-

5:23 And the priest shall write these curses in a book, and mur et respondebit mulier amen amen 5:23 scribetque sahe shall blot them out with the bitter water:

cerdos in libello ista maledicta et delebit ea aquis amarissi-

5:24 And he shall cause the woman to drink the bitter wa- mis in quas maledicta congessit 5:24 et dabit ei bibere quas

ter that causeth the curse: and the water that causeth the curse shall enter into her, and become bitter.

cum exhauserit 5:25 tollet sacerdos de manu eius sacrifi-

5:25 Then the priest shall take the jealousy offering out of the woman’s hand, and shall wave the offering before the LORD, and offer it upon the altar:

cium zelotypiae et eleuabit illud coram Domino inponetque

5:26 And the priest shall take an handful of the offering, super altare ita dumtaxat ut prius 5:26 pugillum sacrificii

even the memorial thereof, and burn it upon the altar, and afterward shall cause the woman to drink the water.

tollat de eo quod offertur et incendat super altare et sic po-

5:27 And when he hath made her to drink the water, then tum det mulieri aquas amarissimas 5:27 quas cum biberit

it shall come to pass, that, if she be defiled, and have done trespass against her husband, that the water that causeth the curse shall enter into her, and become bitter, and her belly shall swell, and her thigh shall rot: and the woman shall be a curse among her people.

si polluta est et contempto uiro adulterii rea pertransibunt eam aquae maledictionis et inflato uentre conputrescet femur eritque mulier in maledictionem et in exemplum omni

5:28 And if the woman be not defiled, but be clean; then populo 5:28 quod si polluta non fuerit erit innoxia et faciet

she shall be free, and shall conceive seed.

5:29 This is the law of jealousies, when a wife goeth aside liberos 5:29 ista est lex zelotypiae si declinauerit mulier a

to another instead of her husband, and is defiled;

5:30 Or when the spirit of jealousy cometh upon him, and uiro suo et si polluta fuerit 5:30 maritusque zelotypiae spi-

he be jealous over his wife, and shall set the woman before the LORD, and the priest shall execute upon her all this

law. ritu concitatus adduxerit eam in conspectu Domini et fecerit

5:31 Then shall the man be guiltless from iniquity, and this ei sacerdos iuxta omnia quae scripta sunt 5:31 maritus abswoman shall bear her iniquity.

que culpa erit et illa recipiet iniquitatem suam

6:1 And the LORD spake unto Moses, saying, 6:1 locutus est Dominus ad Mosen dicens 6:2 loquere ad

6:2 Speak unto the children of Israel, and say unto them, When either man or woman shall separate themselves to vow a vow of a Nazarite, to separate themselves unto the LORD:

filios Israhel et dices ad eos uir siue mulier cum fecerit uo-

6:3 He shall separate himself from wine and strong drink, tum ut sanctificentur et se uoluerint Domino consecrare 6:3

and shall drink no vinegar of wine, or vinegar of strong drink, neither shall he drink any liquor of grapes, nor eat

moist grapes, or dried. uino et omni quod inebriare potest abstinebunt acetum ex uino et ex qualibet alia potione et quicquid de uua exprimitur

6:4 All the days of his separation shall he eat nothing that non bibent uuas recentes siccasque non comedent 6:4 cuncis made of the vine tree, from the kernels even to the husk.

tis diebus quibus ex uoto Domino consecrantur quicquid ex uinea esse potest ab uua passa usque ad acinum non co-

medent 6:5 omni tempore separationis suae nouacula non 6:5 All the days of the vow of his separation there shall no razor come upon his head: until the days be fulfilled, in the which he separateth himself unto the LORD, he shall

transibit super caput eius usque ad conpletum diem quo Do- be holy, and shall let the locks of the hair of his head grow. mino consecratur sanctus erit crescente caesarie capitis eius 6:6 omni tempore consecrationis suae super mortuum non in- 6:6 All the days that he separateth himself unto the LORD

he shall come at no dead body.

gredietur 6:7 nec super patris quidem et matris et fratris so- 6:7 He shall not make himself unclean for his father, or for

his mother, for his brother, or for his sister, when they die: because the consecration of his God is upon his head.

rorisque funere contaminabitur quia consecratio Dei sui super caput eius est 6:8 omnes dies separationis suae sanctus 6:8 All the days of his separation he is holy unto the

LORD.

erit Domino 6:9 sin autem mortuus fuerit subito quispiam 6:9 And if any man die very suddenly by him, and he hath

defiled the head of his consecration; then he shall shave his head in the day of his cleansing, on the seventh day

coram eo polluetur caput consecrationis eius quod radet il- shall he shave it. ico et in eadem die purgationis suae et rursum septima 6:10 6:10 And on the eighth day he shall bring two turtles, or

two young pigeons, to the priest, to the door of the tabernacle of the congregation:

in octauo autem die offeret duos turtures uel duos pullos columbae sacerdoti in introitu foederis testimonii 6:11 faciet- 6:11 And the priest shall offer the one for a sin offering,

and the other for a burnt offering, and make an atonement for him, for that he sinned by the dead, and shall hallow

que sacerdos unum pro peccato et alterum in holocaustum et his head that same day. deprecabitur pro eo quia peccauit super mortuo sanctificabitque caput eius in die illo 6:12 et consecrabit Domino dies 6:12 And he shall consecrate unto the LORD the days of

his separation, and shall bring a lamb of the first year for a trespass offering: but the days that were before shall be

separationis illius offerens agnum anniculum pro peccato ita lost, because his separation was defiled. tamen ut dies priores irriti fiant quoniam polluta est sanctificatio eius 6:13 ista est lex consecrationis cum dies quos 6:13 And this is the law of the Nazarite, when the days of

his separation are fulfilled: he shall be brought unto the door of the tabernacle of the congregation:

ex uoto decreuerat conplebuntur adducet eum ad ostium tabernaculi foederis 6:14 et offeret oblationem eius Domino 6:14 And he shall offer his offering unto the LORD, one he

lamb of the first year without blemish for a burnt offering, and one ewe lamb of the first year without blemish for a sin

agnum anniculum inmaculatum in holocaustum et ouem an- offering, and one ram without blemish for peace offerings,

niculam inmaculatam pro peccato et arietem inmaculatum

hostiam pacificam 6:15 canistrum quoque panum azymo- 6:15 And a basket of unleavened bread, cakes of fine flour mingled with oil, and wafers of unleavened bread anointed with oil, and their meat offering, and their drink offerings.

rum qui conspersi sunt oleo et lagana absque fermento uncta oleo ac libamina singulorum 6:16 quae offeret sacerdos co- 6:16 And the priest shall bring them before the LORD, and

shall offer his sin offering, and his burnt offering:

ram Domino et faciet tam pro peccato quam in holocaustum 6:17 arietem uero immolabit hostiam pacificam Domino of- 6:17 And he shall offer the ram for a sacrifice of peace

offerings unto the LORD, with the basket of unleavened bread: the priest shall offer also his meat offering, and his

ferens simul canistrum azymorum et libamenta quae ex more drink offering. debentur 6:18 tunc radetur nazareus ante ostium taberna- 6:18 And the Nazarite shall shave the head of his sepa-

ration at the door of the tabernacle of the congregation, and shall take the hair of the head of his separation, and put it in the fire which is under the sacrifice of the peace offerings.

culi foederis caesarie consecrationis suae tolletque capillos eius et ponet super ignem qui est subpositus sacrificio pacifi-

Biblia Sacra Vulgata

6:19 And the priest shall take the sodden shoulder of the corum 6:19 et armum coctum arietis tortamque absque fer-

ram, and one unleavened cake out of the basket, and one unleavened wafer, and shall put them upon the hands of the Nazarite, after the hair of his separation is shaven: mento unam de canistro et laganum azymum unum et tradet

6:20 And the priest shall wave them for a wave offering in manibus nazarei postquam rasum fuerit caput eius 6:20

before the LORD: this is holy for the priest, with the wave breast and heave shoulder: and after that the Nazarite may

drink wine. susceptaque rursum ab eo eleuabit in conspectu Domini et sanctificata sacerdotis erunt sicut pectusculum quod separari iussum est et femur post haec potest bibere nazareus uinum

6:21 This is the law of the Nazarite who hath vowed, and 6:21 ista est lex nazarei cum uouerit oblationem suam Do-

of his offering unto the LORD for his separation, beside that that his hand shall get: according to the vow which he

vowed, so he must do after the law of his separation. mino tempore consecrationis suae exceptis his quae inuene-

rit manus eius iuxta quod mente deuouerat ita faciet ad per-

6:22 And the LORD spake unto Moses, saying, fectionem sanctificationis suae 6:22 locutus est Dominus

6:23 Speak unto Aaron and unto his sons, saying, On this ad Mosen dicens 6:23 loquere Aaron et filiis eius sic bewise ye shall bless the children of Israel, saying unto them,

6:24 The LORD bless thee, and keep thee: nedicetis filiis Israhel et dicetis eis 6:24 benedicat tibi Do-6:25 The LORD make his face shine upon thee, and be minus et custodiat te 6:25 ostendat Dominus faciem suam

gracious unto thee:

6:26 The LORD lift up his countenance upon thee, and tibi et misereatur tui 6:26 conuertat Dominus uultum suum

give thee peace.

6:27 And they shall put my name upon the children of ad te et det tibi pacem 6:27 inuocabunt nomen meum super

Israel, and I will bless them.

filios Israhel et ego benedicam eis

7:1 And it came to pass on the day that Moses had fully 7:1 factum est autem in die qua conpleuit Moses taberna-

set up the tabernacle, and had anointed it, and sanctified it, and all the instruments thereof, both the altar and all the vessels thereof, and had anointed them, and sanctified them;

culum et erexit illud unxitque et sanctificauit cum omnibus

7:2 That the princes of Israel, heads of the house of their uasis suis altare similiter et uasa eius 7:2 obtulerunt prin-

fathers, who were the princes of the tribes, and were over them that were numbered, offered:

cipes Israhel et capita familiarum qui erant per singulas tri-

7:3 And they brought their offering before the LORD, six bus praefecti eorum qui numerati fuerant 7:3 munera co-

covered wagons, and twelve oxen; a wagon for two of the princes, and for each one an ox: and they brought them

before the tabernacle. ram Domino sex plaustra tecta cum duodecim bubus unum

plaustrum obtulere duo duces et unum bouem singuli ob-

7:4 And the LORD spake unto Moses, saying, tuleruntque ea in conspectu tabernaculi 7:4 ait autem Do-7:5 Take it of them, that they may be to do the service minus ad Mosen 7:5 suscipe ab eis ut seruiant in minis-

of the tabernacle of the congregation; and thou shalt give them unto the Levites, to every man according to his ser-

vice. terio tabernaculi et tradas ea Leuitis iuxta ordinem minis-7:6 And Moses took the wagons and the oxen, and gave terii sui 7:6 itaque cum suscepisset Moses plaustra et bo-

them unto the Levites.

7:7 Two wagons and four oxen he gave unto the sons of ues tradidit eos Leuitis 7:7 duo plaustra et quattuor boues

Gershon, according to their service:

7:8 And four wagons and eight oxen he gave unto the sons dedit filiis Gerson iuxta id quod habebant necessarium 7:8

of Merari, according unto their service, under the hand of Ithamar the son of Aaron the priest.

quattuor alia plaustra et octo boues dedit filiis Merari secundum officia et cultum suum sub manu Ithamar filii Aaron sa-

7:9 But unto the sons of Kohath he gave none: because cerdotis 7:9 filiis autem Caath non dedit plaustra et boues

the service of the sanctuary belonging unto them was that they should bear upon their shoulders.

quia in sanctuario seruiunt et onera propriis portant umeris 7:10 igitur obtulerunt duces in dedicationem altaris die 7:10 And the princes offered for dedicating of the altar in

the day that it was anointed, even the princes offered their offering before the altar.

qua unctum est oblationem suam ante altare 7:11 dixitque

7:11 And the LORD said unto Moses, They shall offer their offering, each prince on his day, for the dedicating of the altar.

Dominus ad Mosen singuli duces per singulos dies offerant munera in dedicationem altaris 7:12 primo die obtulit ob- 7:12 And he that offered his offering the first day was

Nahshon the son of Amminadab, of the tribe of Judah:

lationem suam Naasson filius Aminadab de tribu Iuda 7:13 7:13 And his offering was one silver charger, the weight

thereof was an hundred and thirty shekels, one silver bowl of seventy shekels, after the shekel of the sanctuary; both of them were full of fine flour mingled with oil for a meat offering:

fueruntque in ea acetabulum argenteum pondo centum triginta siclorum fiala argentea habens septuaginta siclos iuxta pondus sanctuarii utrumque plenum simila conspersa oleo in sacrificium 7:14 mortariolum ex decem siclis aureis ple- 7:14 One spoon of ten shekels of gold, full of incense: num incenso 7:15 bouem et arietem et agnum anniculum in 7:15 One young bullock, one ram, one lamb of the first

year, for a burnt offering:

holocaustum 7:16 hircumque pro peccato 7:17 et in sac- 7:16 One kid of the goats for a sin offering:

7:17 And for a sacrifice of peace offerings, two oxen, five rams, five he goats, five lambs of the first year: this was the offering of Nahshon the son of Amminadab.

rificio pacificorum boues duos arietes quinque hircos quinque agnos anniculos quinque haec est oblatio Naasson filii Aminadab 7:18 secundo die obtulit Nathanahel filius Suar 7:18 On the second day Nethaneel the son of Zuar, prince

of Issachar, did offer:

dux de tribu Isachar 7:19 acetabulum argenteum adpendens 7:19 He offered for his offering one silver charger, the

weight whereof was an hundred and thirty shekels, one silver bowl of seventy shekels, after the shekel of the sanctuary; both of them full of fine flour mingled with oil for a meat offering:

centum triginta siclos fialam argenteam habentem septuaginta siclos iuxta pondus sanctuarii utrumque plenum simila conspersa oleo in sacrificium 7:20 mortariolum aureum ha- 7:20 One spoon of gold of ten shekels, full of incense: bens decem siclos plenum incenso 7:21 bouem de armento 7:21 One young bullock, one ram, one lamb of the first

year, for a burnt offering:

et arietem et agnum anniculum in holocaustum 7:22 hir- 7:22 One kid of the goats for a sin offering: cumque pro peccato 7:23 et in sacrificio pacificorum boues 7:23 And for a sacrifice of peace offerings, two oxen, five

rams, five he goats, five lambs of the first year: this was the offering of Nethaneel the son of Zuar.

duos arietes quinque hircos quinque agnos anniculos quinque haec fuit oblatio Nathanahel filii Suar 7:24 tertio die 7:24 On the third day Eliab the son of Helon, prince of the

children of Zebulun, did offer:

princeps filiorum Zabulon Heliab filius Helon 7:25 obtulit 7:25 His offering was one silver charger, the weight

whereof was an hundred and thirty shekels, one silver bowl of seventy shekels, after the shekel of the sanctuary; both of them full of fine flour mingled with oil for a meat offering:

acetabulum argenteum adpendens centum triginta siclos fialam argenteam habentem septuaginta siclos ad pondus sanctuarii utrumque plenum simila conspersa oleo in sacrificium 7:26 mortariolum aureum adpendens decem siclos plenum 7:26 One golden spoon of ten shekels, full of incense: incenso 7:27 bouem de armento et arietem et agnum anni- 7:27 One young bullock, one ram, one lamb of the first

year, for a burnt offering:

culum in holocaustum 7:28 hircumque pro peccato 7:29 et 7:28 One kid of the goats for a sin offering:

7:29 And for a sacrifice of peace offerings, two oxen, five rams, five he goats, five lambs of the first year: this was the offering of Eliab the son of Helon.

in sacrificio pacificorum boues duos arietes quinque hircos quinque agnos anniculos quinque haec est oblatio Heliab filii Helon 7:30 die quarto princeps filiorum Ruben Helisur 7:30 On the fourth day Elizur the son of Shedeur, prince

of the children of Reuben, did offer:

Biblia Sacra Vulgata

7:31 His offering was one silver charger of the weight of filius Sedeur 7:31 obtulit acetabulum argenteum adpendens

an hundred and thirty shekels, one silver bowl of seventy shekels, after the shekel of the sanctuary; both of them full

of fine flour mingled with oil for a meat offering: centum triginta siclos fialam argenteam habentem septua-

ginta siclos ad pondus sanctuarii utrumque plenum simila

7:32 One golden spoon of ten shekels, full of incense: conspersa oleo in sacrificium 7:32 mortariolum aureum ad-7:33 One young bullock, one ram, one lamb of the first pendens decem siclos plenum incenso 7:33 bouem de ar-

year, for a burnt offering:

7:34 One kid of the goats for a sin offering: mento et arietem et agnum anniculum in holocaustum 7:34 7:35 And for a sacrifice of peace offerings, two oxen, five hircumque pro peccato 7:35 et in hostias pacificorum boues

rams, five he goats, five lambs of the first year: this was the offering of Elizur the son of Shedeur.

duos arietes quinque hircos quinque agnos anniculos quin-

7:36 On the fifth day Shelumiel the son of Zurishaddai, que haec fuit oblatio Helisur filii Sedeur 7:36 die quinto

prince of the children of Simeon, did offer:

7:37 His offering was one silver charger, the weight princeps filiorum Symeon Salamihel filius Surisaddai 7:37

whereof was an hundred and thirty shekels, one silver bowl of seventy shekels, after the shekel of the sanctuary; both of them full of fine flour mingled with oil for a meat offering:

obtulit acetabulum argenteum adpendens centum triginta siclos fialam argenteam habentem septuaginta siclos ad pondus sanctuarii utrumque plenum simila conspersa oleo in sa-

7:38 One golden spoon of ten shekels, full of incense: crificium 7:38 mortariolum aureum adpendens decem sic-7:39 One young bullock, one ram, one lamb of the first los plenum incenso 7:39 bouem de armento et arietem et

year, for a burnt offering:

7:40 One kid of the goats for a sin offering: agnum anniculum in holocaustum 7:40 hircumque pro pec-7:41 And for a sacrifice of peace offerings, two oxen, five cato 7:41 et in hostias pacificorum boues duos arietes quin-

rams, five he goats, five lambs of the first year: this was the offering of Shelumiel the son of Zurishaddai.

que hircos quinque agnos anniculos quinque haec fuit obla-

7:42 On the sixth day Eliasaph the son of Deuel, prince of tio Salamihel filii Surisaddai 7:42 die sexto princeps filio-

the children of Gad, offered:

7:43 His offering was one silver charger of the weight of rum Gad Heliasaph filius Duhel 7:43 obtulit acetabulum ar-

an hundred and thirty shekels, a silver bowl of seventy shekels, after the shekel of the sanctuary; both of them

full of fine flour mingled with oil for a meat offering: genteum adpendens centum triginta siclos fialam argenteam

habentem septuaginta siclos ad pondus sanctuarii utrumque

7:44 One golden spoon of ten shekels, full of incense: plenum simila conspersa oleo in sacrificium 7:44 mortario-7:45 One young bullock, one ram, one lamb of the first lum aureum adpendens siclos decem plenum incenso 7:45

year, for a burnt offering:

bouem de armento et arietem et agnum anniculum in ho-

7:46 One kid of the goats for a sin offering: locaustum 7:46 hircumque pro peccato 7:47 et in hostias

7:47 And for a sacrifice of peace offerings, two oxen, five rams, five he goats, five lambs of the first year: this was the offering of Eliasaph the son of Deuel.

pacificorum boues duos arietes quinque hircos quinque agnos anniculos quinque haec fuit oblatio Heliasaph filii Du-

7:48 On the seventh day Elishama the son of Ammihud, hel 7:48 die septimo princeps filiorum Ephraim Helisama

prince of the children of Ephraim, offered:

7:49 His offering was one silver charger, the weight filius Ammiud 7:49 obtulit acetabulum argenteum adpen-

whereof was an hundred and thirty shekels, one silver bowl of seventy shekels, after the shekel of the sanctuary; both of them full of fine flour mingled with oil for a meat offering:

dens centum triginta siclos fialam argenteam habentem septuaginta siclos ad pondus sanctuarii utrumque plenum simila

7:50 One golden spoon of ten shekels, full of incense: conspersa oleo in sacrificium 7:50 mortariolum aureum adpendens decem siclos plenum incenso 7:51 bouem de ar- 7:51 One young bullock, one ram, one lamb of the first

year, for a burnt offering:

mento et arietem et agnum anniculum in holocaustum 7:52 7:52 One kid of the goats for a sin offering: hircumque pro peccato 7:53 et in hostias pacificas boues 7:53 And for a sacrifice of peace offerings, two oxen, five

rams, five he goats, five lambs of the first year: this was the offering of Elishama the son of Ammihud.

duos arietes quinque hircos quinque agnos anniculos quinque haec fuit oblatio Helisama filii Ammiud 7:54 die oc- 7:54 On the eighth day offered Gamaliel the son of

Pedahzur, prince of the children of Manasseh:

tauo princeps filiorum Manasse Gamalihel filius Phadassur 7:55 obtulit acetabulum argenteum adpendens centum tri- 7:55 His offering was one silver charger of the weight of

an hundred and thirty shekels, one silver bowl of seventy shekels, after the shekel of the sanctuary; both of them full

ginta siclos fialam argenteam habentem septuaginta siclos ad of fine flour mingled with oil for a meat offering: pondus sanctuarii utrumque plenum simila conspersa oleo in sacrificium 7:56 mortariolum aureum adpendens decem 7:56 One golden spoon of ten shekels, full of incense: siclos plenum incenso 7:57 bouem de armento et arietem 7:57 One young bullock, one ram, one lamb of the first

year, for a burnt offering:

et agnum anniculum in holocaustum 7:58 hircumque pro 7:58 One kid of the goats for a sin offering: peccato 7:59 et in hostias pacificorum boues duos arietes 7:59 And for a sacrifice of peace offerings, two oxen, five

rams, five he goats, five lambs of the first year: this was the offering of Gamaliel the son of Pedahzur.

quinque hircos quinque agnos anniculos quinque haec fuit oblatio Gamalihel filii Phadassur 7:60 die nono princeps fi- 7:60 On the ninth day Abidan the son of Gideoni, prince

of the children of Benjamin, offered:

liorum Beniamin Abidan filius Gedeonis 7:61 obtulit ace- 7:61 His offering was one silver charger, the weight

whereof was an hundred and thirty shekels, one silver bowl of seventy shekels, after the shekel of the sanctuary; both of them full of fine flour mingled with oil for a meat offering:

tabulum argenteum adpendens centum triginta siclos fialam argenteam habentem septuaginta siclos ad pondus sanctuarii utrumque plenum simila conspersa oleo in sacrificium 7:62 mortariolum aureum adpendens decem siclos plenum 7:62 One golden spoon of ten shekels, full of incense: incenso 7:63 bouem de armento et arietem et agnum an- 7:63 One young bullock, one ram, one lamb of the first

year, for a burnt offering:

niculum in holocaustum 7:64 hircumque pro peccato 7:65 7:64 One kid of the goats for a sin offering:

7:65 And for a sacrifice of peace offerings, two oxen, five rams, five he goats, five lambs of the first year: this was the offering of Abidan the son of Gideoni.

et in hostias pacificorum boues duos arietes quinque hircos quinque agnos anniculos quinque haec fuit oblatio Abidan filii Gedeonis 7:66 die decimo princeps filiorum Dan Ahie- 7:66 On the tenth day Ahiezer the son of Ammishaddai,

prince of the children of Dan, offered:

zer filius Amisaddai 7:67 obtulit acetabulum argenteum ad- 7:67 His offering was one silver charger, the weight

whereof was an hundred and thirty shekels, one silver bowl of seventy shekels, after the shekel of the sanctuary; both of them full of fine flour mingled with oil for a meat offering:

pendens centum triginta siclos fialam argenteam habentem septuaginta siclos ad pondus sanctuarii utrumque plenum simila conspersa oleo in sacrificium 7:68 mortariolum au- 7:68 One golden spoon of ten shekels, full of incense: reum adpendens decem siclos plenum incenso 7:69 bouem 7:69 One young bullock, one ram, one lamb of the first

year, for a burnt offering:

de armento et arietem et agnum anniculum in holocaustum 7:70 hircumque pro peccato 7:71 et in hostias pacificorum 7:70 One kid of the goats for a sin offering:

7:71 And for a sacrifice of peace offerings, two oxen, five rams, five he goats, five lambs of the first year: this was the offering of Ahiezer the son of Ammishaddai.

boues duos arietes quinque hircos quinque agnos anniculos

Biblia Sacra Vulgata

quinque haec fuit oblatio Ahiezer filii Amisaddai

7:72 On the eleventh day Pagiel the son of Ocran, prince 7:72 die undecimo princeps filiorum Aser Phagaihel filius

of the children of Asher, offered:

7:73 His offering was one silver charger, the weight Ochran 7:73 obtulit acetabulum argenteum adpendens cen-

whereof was an hundred and thirty shekels, one silver bowl of seventy shekels, after the shekel of the sanctuary; both of them full of fine flour mingled with oil for a meat offering:

tum triginta siclos fialam argenteam habentem septuaginta siclos ad pondus sanctuarii utrumque plenum simila con-

7:74 One golden spoon of ten shekels, full of incense: spersa oleo in sacrificium 7:74 mortariolum aureum adpen-7:75 One young bullock, one ram, one lamb of the first dens decem siclos plenum incenso 7:75 bouem de armento

year, for a burnt offering:

7:76 One kid of the goats for a sin offering: et arietem et agnum anniculum in holocaustum 7:76 hir-

7:77 And for a sacrifice of peace offerings, two oxen, five cumque pro peccato 7:77 et in hostias pacificorum boues rams, five he goats, five lambs of the first year: this was the offering of Pagiel the son of Ocran.

duos arietes quinque hircos quinque agnos anniculos quin-

7:78 On the twelfth day Ahira the son of Enan, prince of que haec fuit oblatio Phagaihel filii Ochran 7:78 die duode-

the children of Naphtali, offered:

7:79 His offering was one silver charger, the weight cimo princeps filiorum Nepthalim Achira filius Henan 7:79

whereof was an hundred and thirty shekels, one silver bowl of seventy shekels, after the shekel of the sanctuary; both of them full of fine flour mingled with oil for a meat offering:

obtulit acetabulum argenteum adpendens centum triginta siclos fialam argenteam habentem septuaginta siclos ad pondus sanctuarii utrumque plenum simila conspersa oleo in sa-

7:80 One golden spoon of ten shekels, full of incense: crificium 7:80 mortariolum aureum adpendens decem sic-7:81 One young bullock, one ram, one lamb of the first los plenum incenso 7:81 bouem de armento et arietem et

year, for a burnt offering:

7:82 One kid of the goats for a sin offering: agnum anniculum in holocaustum 7:82 hircumque pro pec-7:83 And for a sacrifice of peace offerings, two oxen, five cato 7:83 et in hostias pacificorum boues duos arietes quin-

rams, five he goats, five lambs of the first year: this was the offering of Ahira the son of Enan.

que hircos quinque agnos anniculos quinque haec fuit obla-

7:84 This was the dedication of the altar, in the day when tio Achira filii Henan 7:84 haec in dedicatione altaris oblata

it was anointed, by the princes of Israel: twelve chargers of silver, twelve silver bowls, twelve spoons of gold:

sunt a principibus Israhel in die qua consecratum est acetabula argentea duodecim fialae argenteae duodecim mor-

7:85 Each charger of silver weighing an hundred and thirty tariola aurea duodecim 7:85 ita ut centum triginta argenti

shekels, each bowl seventy: all the silver vessels weighed two thousand and four hundred shekels, after the shekel of

the sanctuary: siclos haberet unum acetabulum et septuaginta siclos una fi-

ala id est in commune uasorum omnium ex argento sicli duo

7:86 The golden spoons were twelve, full of incense, milia quadringenti pondere sanctuarii 7:86 mortariola au-

weighing ten shekels apiece, after the shekel of the sanctuary: all the gold of the spoons was an hundred and twenty

shekels. rea duodecim plena incenso denos siclos adpendentia pon-7:87 All the oxen for the burnt offering were twelve bul- dere sanctuarii id est simul auri sicli centum uiginti 7:87

locks, the rams twelve, the lambs of the first year twelve, with their meat offering: and the kids of the goats for sin

offering twelve. boues de armento in holocaustum duodecim arietes duode-

cim agni anniculi duodecim et libamenta eorum hirci duo-

7:88 And all the oxen for the sacrifice of the peace offer- decim pro peccato 7:88 hostiae pacificorum boues uiginti

ings were twenty and four bullocks, the rams sixty, the he goats sixty, the lambs of the first year sixty. This was the

dedication of the altar, after that it was anointed. quattuor arietes sexaginta hirci sexaginta agni anniculi sexaginta haec oblata sunt in dedicatione altaris quando unctum est 7:89 cumque ingrederetur Moses tabernaculum foede- 7:89 And when Moses was gone into the tabernacle of the

congregation to speak with him, then he heard the voice of one speaking unto him from off the mercy seat that was upon the ark of testimony, from between the two cherubims: and he spake unto him.

ris ut consuleret oraculum audiebat uocem loquentis ad se de propitiatorio quod erat super arcam testimonii inter duos cherubin unde et loquebatur ei

8:1 locutus est Dominus ad Mosen dicens 8:2 loquere 8:1 And the LORD spake unto Moses, saying,

8:2 Speak unto Aaron and say unto him, When thou lightest the lamps, the seven lamps shall give light over against the candlestick.

Aaroni et dices ad eum cum posueris septem lucernas contra eam partem quam candelabrum respicit lucere debebunt 8:3 fecitque Aaron et inposuit lucernas super candelabrum 8:3 And Aaron did so; he lighted the lamps thereof over

against the candlestick, as the LORD commanded Moses.

ut praeceperat Dominus Mosi 8:4 haec autem erat factura 8:4 And this work of the candlestick was of beaten gold,

unto the shaft thereof, unto the flowers thereof, was beaten work: according unto the pattern which the LORD had

candelabri ex auro ductili tam medius stipes quam cuncta ex shewed Moses, so he made the candlestick. utroque calamorum latera nascebantur iuxta exemplum quod ostendit Dominus Mosi ita operatus est candelabrum 8:5 et 8:5 And the LORD spake unto Moses, saying, locutus est Dominus ad Mosen dicens 8:6 tolle Leuitas de 8:6 Take the Levites from among the children of Israel,

and cleanse them.

medio filiorum Israhel et purificabis eos 8:7 iuxta hunc ri- 8:7 And thus shalt thou do unto them, to cleanse them:

Sprinkle water of purifying upon them, and let them shave all their flesh, and let them wash their clothes, and so make

tum aspergantur aqua lustrationis et radant omnes pilos car- themselves clean. nis suae cumque lauerint uestimenta sua et mundati fuerint 8:8 tollant bouem de armentis et libamentum eius similam 8:8 Then let them take a young bullock with his meat of-

fering, even fine flour mingled with oil, and another young bullock shalt thou take for a sin offering.

oleo conspersam bouem autem alterum de armento tu accipies pro peccato 8:9 et adplicabis Leuitas coram taberna- 8:9 And thou shalt bring the Levites before the taberna-

cle of the congregation: and thou shalt gather the whole assembly of the children of Israel together:

culo foederis conuocata omni multitudine filiorum Israhel 8:10 cumque Leuitae fuerint coram Domino ponent filii Isra- 8:10 And thou shalt bring the Levites before the LORD:

and the children of Israel shall put their hands upon the Levites:

hel manus suas super eos 8:11 et offeret Aaron Leuitas mu-

8:11 And Aaron shall offer the Levites before the LORD for an offering of the children of Israel, that they may execute the service of the LORD.

nus in conspectu Domini a filiis Israhel ut seruiant in ministerio eius 8:12 Leuitae quoque ponent manus suas super ca- 8:12 And the Levites shall lay their hands upon the heads

of the bullocks: and thou shalt offer the one for a sin offering, and the other for a burnt offering, unto the LORD,

pita boum e quibus unum facies pro peccato et alterum in ho- to make an atonement for the Levites. locaustum Domini ut depreceris pro eis 8:13 statuesque Le- 8:13 And thou shalt set the Levites before Aaron, and

before his sons, and offer them for an offering unto the LORD.

uitas in conspectu Aaron et filiorum eius et consecrabis oblatos Domino 8:14 ac separabis de medio filiorum Israhel ut 8:14 Thus shalt thou separate the Levites from among the

children of Israel: and the Levites shall be mine.

sint mei 8:15 et postea ingrediantur tabernaculum foederis 8:15 And after that shall the Levites go in to do the ser-

vice of the tabernacle of the congregation: and thou shalt cleanse them, and offer them for an offering.

ut seruiant mihi sicque purificabis et consecrabis eos in oblationem Domini quoniam dono donati sunt mihi a filiis Israhel 8:16 pro primogenitis quae aperiunt omnem uuluam in 8:16 For they are wholly given unto me from among the

children of Israel; instead of such as open every womb, even instead of the firstborn of all the children of Israel, have I taken them unto me.

Biblia Sacra Vulgata

8:17 For all the firstborn of the children of Israel are mine, Israhel accepi eos 8:17 mea sunt omnia primogenita filioboth man and beast: on the day that I smote every firstborn in the land of Egypt I sanctified them for myself.

rum Israhel tam ex hominibus quam ex iumentis ex die quo percussi omnem primogenitum in terra Aegypti sanctificaui

8:18 And I have taken the Levites for all the firstborn of eos mihi 8:18 et tuli Leuitas pro cunctis primogenitis filio-

the children of Israel.

8:19 And I have given the Levites as a gift to Aaron and rum Israhel 8:19 tradidique eos dono Aaroni et filiis eius

to his sons from among the children of Israel, to do the service of the children of Israel in the tabernacle of the congregation, and to make an atonement for the children of Israel: that there be no plague among the children of Israel, when the children of Israel come nigh unto the sanctuary.

de medio populi ut seruiant mihi pro Israhel in tabernaculo foederis et orent pro eis ne sit in populo plaga si ausi fuerint

8:20 And Moses, and Aaron, and all the congregation of accedere ad sanctuarium 8:20 feceruntque Moses et Aaron

the children of Israel, did to the Levites according unto all that the LORD commanded Moses concerning the Levites,

so did the children of Israel unto them. et omnis multitudo filiorum Israhel super Leuitas quae prae-8:21 And the Levites were purified, and they washed their ceperat Dominus Mosi 8:21 purificatique sunt et lauerunt

clothes; and Aaron offered them as an offering before the LORD; and Aaron made an atonement for them to cleanse

them. uestimenta sua eleuauitque eos Aaron in conspectu Domini 8:22 And after that went the Levites in to do their ser- et orauit pro eis 8:22 ut purificati ingrederentur ad officia

vice in the tabernacle of the congregation before Aaron, and before his sons: as the LORD had commanded Moses

concerning the Levites, so did they unto them. sua in tabernaculum foederis coram Aaron et filiis eius sicut 8:23 And the LORD spake unto Moses, saying, praeceperat Dominus Mosi de Leuitis ita factum est 8:23

8:24 This is it that belongeth unto the Levites: from twenty locutus est Dominus ad Mosen dicens 8:24 haec est lex and five years old and upward they shall go in to wait upon the service of the tabernacle of the congregation:

Leuitarum a uiginti quinque annis et supra ingredientur ut

8:25 And from the age of fifty years they shall cease wait- ministrent in tabernaculo foederis 8:25 cumque quinqua-

ing upon the service thereof, and shall serve no more: 8:26 But shall minister with their brethren in the taberna- gesimum annum aetatis impleuerint seruire cessabunt 8:26 cle of the congregation, to keep the charge, and shall do no service. Thus shalt thou do unto the Levites touching

their charge. eruntque ministri fratrum suorum in tabernaculo foederis ut custodiant quae sibi fuerint commendata opera autem ipsa non faciant sic dispones Leuitas in custodiis suis

9:1 And the LORD spake unto Moses in the wilderness of 9:1 locutus est Dominus ad Mosen in deserto Sinai anno

Sinai, in the first month of the second year after they were come out of the land of Egypt, saying,

secundo postquam egressi sunt de terra Aegypti mense

9:2 Let the children of Israel also keep the passover at his primo dicens 9:2 faciant filii Israhel phase in tempore suo appointed season.

9:3 In the fourteenth day of this month, at even, ye shall 9:3 quartadecima die mensis huius ad uesperam iuxta omnes

keep it in his appointed season: according to all the rites of it, and according to all the ceremonies thereof, shall ye

keep it. caerimonias et iustificationes eius 9:4 praecepitque Moses

9:4 And Moses spake unto the children of Israel, that they should keep the passover.

filiis Israhel ut facerent phase 9:5 qui fecerunt tempore suo

9:5 And they kept the passover on the fourteenth day of the first month at even in the wilderness of Sinai: according to all that the LORD commanded Moses, so did the children of Israel.

quartadecima die mensis ad uesperam in monte Sinai iuxta omnia quae mandauerat Dominus Mosi fecerunt filii Israhel

9:6 And there were certain men, who were defiled by the 9:6 ecce autem quidam inmundi super animam hominis qui

dead body of a man, that they could not keep the passover on that day: and they came before Moses and before

Aaron on that day: non poterant facere pascha in die illo accedentes ad Mosen 9:7 And those men said unto him, We are defiled by the et Aaron 9:7 dixerunt eis inmundi sumus super animam ho-

dead body of a man: wherefore are we kept back, that we may not offer an offering of the LORD in his appointed

season among the children of Israel? minis quare fraudamur ut non ualeamus offerre oblationem Domino in tempore suo inter filios Israhel 9:8 quibus res- 9:8 And Moses said unto them, Stand still, and I will hear

what the LORD will command concerning you.

pondit Moses state ut consulam quid praecipiat Dominus de uobis 9:9 locutusque est Dominus ad Mosen dicens 9:10 9:9 And the LORD spake unto Moses, saying,

9:10 Speak unto the children of Israel, saying, If any man of you or of your posterity shall be unclean by reason of a dead body, or be in a journey afar off, yet he shall keep the passover unto the LORD.

loquere filiis Israhel homo qui fuerit inmundus super anima siue in uia procul in gente uestra faciat phase Domino 9:11 _{9:11 The fourteenth day of the second month at even they}

shall keep it, and eat it with unleavened bread and bitter herbs.

mense secundo quartadecima die mensis ad uesperam cum azymis et lactucis agrestibus comedent illud 9:12 non re- 9:12 They shall leave none of it unto the morning, nor

break any bone of it: according to all the ordinances of the passover they shall keep it.

linquent ex eo quippiam usque mane et os eius non confringent omnem ritum phase obseruabunt 9:13 si quis autem et 9:13 But the man that is clean, and is not in a journey, and

forbeareth to keep the passover, even the same soul shall be cut off from among his people: because he brought not the offering of the LORD in his appointed season, that man shall bear his sin.

mundus est et in itinere non fuit et tamen non fecit phase exterminabitur anima illa de populis suis quia sacrificium Domino non obtulit tempore suo peccatum suum ipse portabit 9:14 peregrinus quoque et aduena si fuerit apud uos 9:14 And if a stranger shall sojourn among you, and will

keep the passover unto the LORD; according to the ordinance of the passover, and according to the manner thereof, so shall he do: ye shall have one ordinance, both for the stranger, and for him that was born in the land.

faciet phase Domini iuxta caerimonias et iustificationes eius praeceptum idem erit apud uos tam aduenae quam indigenae 9:15 igitur die qua erectum est tabernaculum operuit illud nu- 9:15 And on the day that the tabernacle was reared up the

cloud covered the tabernacle, namely, the tent of the testimony: and at even there was upon the tabernacle as it were

bes a uespere autem super tentorium erat quasi species ignis the appearance of fire, until the morning. usque mane 9:16 sic fiebat iugiter per diem operiebat illud 9:16 So it was alway: the cloud covered it by day, and the

appearance of fire by night.

nubes et per noctem quasi species ignis 9:17 cumque ablata 9:17 And when the cloud was taken up from the taberna-

cle, then after that the children of Israel journeyed: and in the place where the cloud abode, there the children of

fuisset nubes quae tabernaculum protegebat tunc proficisce- Israel pitched their tents. bantur filii Israhel et in loco ubi stetisset nubes ibi castrametabantur 9:18 ad imperium Domini proficiscebantur et ad 9:18 At the commandment of the LORD the children of

Israel journeyed, and at the commandment of the LORD they pitched: as long as the cloud abode upon the taberna-

imperium illius figebant tabernaculum cunctis diebus quibus cle they rested in their tents. stabat nubes super tabernaculum manebant in eodem loco 9:19 et si euenisset ut multo tempore maneret super illud er- 9:19 And when the cloud tarried long upon the tabernacle

many days, then the children of Israel kept the charge of the LORD, and journeyed not.

ant filii Israhel in excubiis Domini et non proficiscebantur 9:20 quotquot diebus fuisset nubes super tabernaculum ad 9:20 And so it was, when the cloud was a few days

upon the tabernacle; according to the commandment of the LORD they abode in their tents, and according to the

imperium Domini erigebant tentoria et ad imperium illius commandment of the LORD they journeyed. deponebant 9:21 si fuisset nubes a uespere usque mane et 9:21 And so it was, when the cloud abode from even unto

the morning, and that the cloud was taken up in the morning, then they journeyed: whether it was by day or by night

statim diluculo tabernaculum reliquisset proficiscebantur et that the cloud was taken up, they journeyed. si post diem et noctem recessisset dissipabant tentoria 9:22 9:22 Or whether it were two days, or a month, or a

year, that the cloud tarried upon the tabernacle, remaining thereon, the children of Israel abode in their tents, and

si biduo aut uno mense uel longiori tempore fuisset super journeyed not: but when it was taken up, they journeyed. tabernaculum manebant filii Israhel in eodem loco et non

Biblia Sacra Vulgata

proficiscebantur statim autem ut recessisset mouebant castra

9:23 At the commandment of the LORD they rested in the 9:23 per uerbum Domini figebant tentoria et per uerbum il-

tents, and at the commandment of the LORD they journeyed: they kept the charge of the LORD, at the com-

mandment of the LORD by the hand of Moses. lius proficiscebantur erantque in excubiis Domini iuxta imperium eius per manum Mosi

10:1 And the LORD spake unto Moses, saying, 10:1 locutus est Dominus ad Mosen dicens 10:2 fac tibi

10:2 Make thee two trumpets of silver; of a whole piece shalt thou make them: that thou mayest use them for the calling of the assembly, and for the journeying of the camps.

duas tubas argenteas ductiles quibus conuocare possis mul-

10:3 And when they shall blow with them, all the assem- titudinem quando mouenda sunt castra 10:3 cumque inbly shall assemble themselves to thee at the door of the tabernacle of the congregation.

crepueris tubis congregabitur ad te omnis turba ad ostium

10:4 And if they blow but with one trumpet, then the foederis tabernaculi 10:4 si semel clangueris uenient ad te

princes, which are heads of the thousands of Israel, shall gather themselves unto thee.

principes et capita multitudinis Israhel 10:5 sin autem pro-

10:5 When ye blow an alarm, then the camps that lie on the east parts shall go forward.

lixior atque concisus clangor increpuerit mouebunt castra

10:6 When ye blow an alarm the second time, then the primi qui sunt ad orientalem plagam 10:6 in secundo au-

camps that lie on the south side shall take their journey: they shall blow an alarm for their journeys.

tem sonitu et pari ululatu tubae leuabunt tentoria qui habitant ad meridiem et iuxta hunc modum reliqui facient ulu-

10:7 But when the congregation is to be gathered together, lantibus tubis in profectione 10:7 quando autem congregan-

ye shall blow, but ye shall not sound an alarm.

dus est populus simplex tubarum clangor erit et non concise

10:8 And the sons of Aaron, the priests, shall blow with ululabunt 10:8 filii Aaron sacerdotes clangent tubis eritque

the trumpets; and they shall be to you for an ordinance for ever throughout your generations.

hoc legitimum sempiternum in generationibus uestris 10:9

10:9 And if ye go to war in your land against the enemy that oppresseth you, then ye shall blow an alarm with the trumpets; and ye shall be remembered before the LORD your God, and ye shall be saved from your enemies.

si exieritis ad bellum de terra uestra contra hostes qui dimicant aduersum uos clangetis ululantibus tubis et erit recordatio uestri coram Domino Deo uestro ut eruamini de manibus

10:10 Also in the day of your gladness, and in your solemn inimicorum uestrorum 10:10 si quando habebitis epulum et

days, and in the beginnings of your months, ye shall blow with the trumpets over your burnt offerings, and over the sacrifices of your peace offerings; that they may be to you for a memorial before your God: I am the LORD your God.

dies festos et kalendas canetis tubis super holocaustis et pacificis uictimis ut sint uobis in recordationem Dei uestri ego

10:11 And it came to pass on the twentieth day of the sec- Dominus Deus uester 10:11 anno secundo mense secundo ond month, in the second year, that the cloud was taken up from off the tabernacle of the testimony.

uicesima die mensis eleuata est nubes de tabernaculo foe-

10:12 And the children of Israel took their journeys out of deris 10:12 profectique sunt filii Israhel per turmas suas de

the wilderness of Sinai; and the cloud rested in the wilderness of Paran.

deserto Sinai et recubuit nubes in solitudine Pharan 10:13

10:13 And they first took their journey according to the commandment of the LORD by the hand of Moses.

moueruntque castra primi iuxta imperium Domini in manu

10:14 In the first place went the standard of the camp of Mosi 10:14 filii Iuda per turmas suas quorum princeps erat

the children of Judah according to their armies: and over his host was Nahshon the son of Amminadab.

Naasson filius Aminadab 10:15 in tribu filiorum Isachar fuit

10:15 And over the host of the tribe of the children of Issachar was Nethaneel the son of Zuar.

10:16 And over the host of the tribe of the children of Ze- princeps Nathanahel filius Suar 10:16 in tribu Zabulon erat

bulun was Eliab the son of Helon.

10:17 And the tabernacle was taken down; and the sons of princeps Heliab filius Helon 10:17 depositumque est taber-

Gershon and the sons of Merari set forward, bearing the tabernacle.

naculum quod portantes egressi sunt filii Gerson et Merari 10:18 profectique sunt et filii Ruben per turmas et ordinem 10:18 And the standard of the camp of Reuben set forward

according to their armies: and over his host was Elizur the son of Shedeur.

suum quorum princeps erat Helisur filius Sedeur 10:19 in

10:19 And over the host of the tribe of the children of Simeon was Shelumiel the son of Zurishaddai.

tribu autem filiorum Symeon princeps fuit Salamihel filius Surisaddai 10:20 porro in tribu Gad erat princeps Heliasaph 10:20 And over the host of the tribe of the children of Gad

was Eliasaph the son of Deuel.

filius Duhel 10:21 profectique sunt et Caathitae portantes 10:21 And the Kohathites set forward, bearing the sanc-

tuary: and the other did set up the tabernacle against they came.

sanctuarium tamdiu tabernaculum portabatur donec uenirent ad erectionis locum 10:22 mouerunt castra et filii Ephraim 10:22 And the standard of the camp of the children of

Ephraim set forward according to their armies: and over his host was Elishama the son of Ammihud.

per turmas suas in quorum exercitu princeps erat Helisama filius Ammiud 10:23 in tribu autem filiorum Manasse prin- 10:23 And over the host of the tribe of the children of

Manasseh was Gamaliel the son of Pedahzur.

ceps fuit Gamalihel filius Phadassur 10:24 et in tribu Benia- 10:24 And over the host of the tribe of the children of Ben-

jamin was Abidan the son of Gideoni.

min dux Abidan filius Gedeonis 10:25 nouissimi castrorum 10:25 And the standard of the camp of the children of

Dan set forward, which was the rereward of all the camps throughout their hosts: and over his host was Ahiezer the

omnium profecti sunt filii Dan per turmas suas in quorum ex- son of Ammishaddai. ercitu princeps fuit Ahiezer filius Amisaddai 10:26 in tribu 10:26 And over the host of the tribe of the children of

Asher was Pagiel the son of Ocran.

autem filiorum Aser erat princeps Phagaihel filius Ochran 10:27 et in tribu filiorum Nepthalim princeps Achira filius 10:27 And over the host of the tribe of the children of

Naphtali was Ahira the son of Enan.

Henan 10:28 haec sunt castra et profectiones filiorum Is- 10:28 Thus were the journeyings of the children of Israel

according to their armies, when they set forward.

rahel per turmas suas quando egrediebantur 10:29 dixitque 10:29 And Moses said unto Hobab, the son of Raguel the

Midianite, Moses’ father in law, We are journeying unto the place of which the LORD said, I will give it you: come thou with us, and we will do thee good: for the LORD hath spoken good concerning Israel.

Moses Hobab filio Rahuhel Madianiti cognato suo proficiscimur ad locum quem Dominus daturus est nobis ueni nobiscum ut benefaciamus tibi quia Dominus bona promisit Israheli 10:30 cui ille respondit non uadam tecum sed re- 10:30 And he said unto him, I will not go; but I will depart

to mine own land, and to my kindred.

uertar in terram meam in qua natus sum 10:31 et ille noli 10:31 And he said, Leave us not, I pray thee; forasmuch

as thou knowest how we are to encamp in the wilderness, and thou mayest be to us instead of eyes.

inquit nos relinquere tu enim nosti in quibus locis per desertum castra ponere debeamus et eris ductor noster 10:32 10:32 And it shall be, if thou go with us, yea, it shall be,

that what goodness the LORD shall do unto us, the same will we do unto thee.

cumque nobiscum ueneris quicquid optimum fuerit ex opibus quas nobis traditurus est Dominus dabimus tibi 10:33 10:33 And they departed from the mount of the LORD

three days’ journey: and the ark of the covenant of the LORD went before them in the three days’ journey, to

profecti sunt ergo de monte Domini uia trium dierum ar- search out a resting place for them. caque foederis Domini praecedebat eos per dies tres prouidens castrorum locum 10:34 nubes quoque Domini super 10:34 And the cloud of the LORD was upon them by day,

when they went out of the camp.

eos erat per diem cum incederent 10:35 cumque eleuaretur 10:35 And it came to pass, when the ark set forward, that

Moses said, Rise up, LORD, and let thine enemies be scattered; and let them that hate thee flee before thee.

arca dicebat Moses surge Domine et dissipentur inimici tui et fugiant qui oderunt te a facie tua 10:36 cum autem depo- 10:36 And when it rested, he said, Return, O LORD, unto

the many thousands of Israel.

Biblia Sacra Vulgata

neretur aiebat reuertere Domine ad multitudinem exercitus Israhel

11:1 And when the people complained, it displeased the 11:1 interea ortum est murmur populi quasi dolentium pro

LORD: and the LORD heard it; and his anger was kindled; and the fire of the LORD burnt among them, and consumed them that were in the uttermost parts of the camp.

labore contra Dominum quod cum audisset iratus est et accensus in eos ignis Domini deuorauit extremam castrorum

11:2 And the people cried unto Moses; and when Moses partem 11:2 cumque clamasset populus ad Mosen orauit

prayed unto the LORD, the fire was quenched. 11:3 And he called the name of the place Taberah: because Moses Dominum et absortus est ignis 11:3 uocauitque nothe fire of the LORD burnt among them.

men loci illius Incensio eo quod succensus fuisset contra

11:4 And the mixt multitude that was among them fell a eos ignis Domini 11:4 uulgus quippe promiscuum quod as-

lusting: and the children of Israel also wept again, and said, Who shall give us flesh to eat?

cenderat cum eis flagrauit desiderio sedens et flens iunctis sibi pariter filiis Israhel et ait quis dabit nobis ad uescendum

11:5 We remember the fish, which we did eat in Egypt carnes 11:5 recordamur piscium quos comedebamus in Ae-

freely; the cucumbers, and the melons, and the leeks, and the onions, and the garlick:

gypto gratis in mentem nobis ueniunt cucumeres et pepones

11:6 But now our soul is dried away: there is nothing at porrique et cepae et alia 11:6 anima nostra arida est nihil

all, beside this manna, before our eyes.

11:7 And the manna was as coriander seed, and the colour aliud respiciunt oculi nostri nisi man 11:7 erat autem man

thereof as the colour of bdellium.

11:8 And the people went about, and gathered it, and quasi semen coriandri coloris bdellii 11:8 circuibatque po-

ground it in mills, or beat it in a mortar, and baked it in pans, and made cakes of it: and the taste of it was as the

taste of fresh oil. pulus et colligens illud frangebat mola siue terebat in mor-

tario coquens in olla et faciens ex eo tortulas saporis quasi

11:9 And when the dew fell upon the camp in the night, panis oleati 11:9 cumque descenderet nocte super castra ros

the manna fell upon it.

11:10 Then Moses heard the people weep throughout their descendebat pariter et man 11:10 audiuit ergo Moses flen-

families, every man in the door of his tent: and the anger of the LORD was kindled greatly; Moses also was dis-

pleased. tem populum per familias singulos per ostia tentorii sui ira-

tusque est furor Domini ualde sed et Mosi intoleranda res

11:11 And Moses said unto the LORD, Wherefore hast uisa est 11:11 et ait ad Dominum cur adflixisti seruum tuum

thou afflicted thy servant? and wherefore have I not found favour in thy sight, that thou layest the burden of all this

people upon me? quare non inuenio gratiam coram te et cur inposuisti pondus 11:12 Have I conceived all this people? have I begotten uniuersi populi huius super me 11:12 numquid ego concepi

them, that thou shouldest say unto me, Carry them in thy bosom, as a nursing father beareth the sucking child, unto

the land which thou swarest unto their fathers? omnem hanc multitudinem uel genui eam ut dicas mihi porta

eos in sinu tuo sicut portare solet nutrix infantulum et defer

11:13 Whence should I have flesh to give unto all this peo- in terram pro qua iurasti patribus eorum 11:13 unde mihi ple? for they weep unto me, saying, Give us flesh, that we may eat.

carnes ut dem tantae multitudini flent contra me dicentes da

11:14 I am not able to bear all this people alone, because nobis carnes ut comedamus 11:14 non possum solus sustiit is too heavy for me. 11:15 And if thou deal thus with me, kill me, I pray thee, nere omnem hunc populum quia grauis mihi est 11:15 sin out of hand, if I have found favour in thy sight; and let me not see my wretchedness.

aliter tibi uidetur obsecro ut interficias me et inueniam gra-

11:16 And the LORD said unto Moses, Gather unto me tiam in oculis tuis ne tantis adficiar malis 11:16 et dixit Do-

seventy men of the elders of Israel, whom thou knowest to be the elders of the people, and officers over them; and bring them unto the tabernacle of the congregation, that they may stand there with thee.

minus ad Mosen congrega mihi septuaginta uiros de senibus Israhel quos tu nosti quod senes populi sint ac magistri et duces eos ad ostium tabernaculi foederis faciesque ibi stare tecum 11:17 ut descendam et loquar tibi et auferam de spi- 11:17 And I will come down and talk with thee there: and

I will take of the spirit which is upon thee, and will put it upon them; and they shall bear the burden of the people

ritu tuo tradamque eis ut sustentent tecum onus populi et non with thee, that thou bear it not thyself alone. tu solus graueris 11:18 populo quoque dices sanctificamini 11:18 And say thou unto the people, Sanctify yourselves

against to morrow, and ye shall eat flesh: for ye have wept in the ears of the LORD, saying, Who shall give us flesh to eat? for it was well with us in Egypt: therefore the LORD will give you flesh, and ye shall eat.

cras comedetis carnes ego enim audiui uos dicere quis dabit nobis escas carnium bene nobis erat in Aegypto ut det uobis Dominus carnes et comedatis 11:19 non uno die nec duobus 11:19 Ye shall not eat one day, nor two days, nor five days,

neither ten days, nor twenty days;

uel quinque aut decem nec uiginti quidem 11:20 sed usque 11:20 But even a whole month, until it come out at your

nostrils, and it be loathsome unto you: because that ye have despised the LORD which is among you, and have wept before him, saying, Why came we forth out of Egypt?

ad mensem dierum donec exeat per nares uestras et uertatur in nausiam eo quod reppuleritis Dominum qui in medio uestri est et fleueritis coram eo dicentes quare egressi sumus ex Aegypto 11:21 et ait Moses sescenta milia peditum 11:21 And Moses said, The people, among whom I am,

are six hundred thousand footmen; and thou hast said, I will give them flesh, that they may eat a whole month.

huius populi sunt et tu dicis dabo eis esum carnium mense integro 11:22 numquid ouium et boum multitudo caedetur 11:22 Shall the flocks and the herds be slain for them, to

suffice them? or shall all the fish of the sea be gathered together for them, to suffice them?

ut possit sufficere ad cibum uel omnes pisces maris in unum congregabuntur ut eos satient 11:23 cui respondit Dominus 11:23 And the LORD said unto Moses, Is the LORD’s

hand waxed short? thou shalt see now whether my word shall come to pass unto thee or not.

numquid manus Domini inualida est iam nunc uidebis utrum meus sermo opere conpleatur 11:24 uenit igitur Moses et 11:24 And Moses went out, and told the people the words

of the LORD, and gathered the seventy men of the elders of the people, and set them round about the tabernacle.

narrauit populo uerba Domini congregans septuaginta uiros de senibus Israhel quos stare fecit circa tabernaculum 11:25 11:25 And the LORD came down in a cloud, and spake

unto him, and took of the spirit that was upon him, and gave it unto the seventy elders: and it came to pass, that, when the spirit rested upon them, they prophesied, and did not cease.

descenditque Dominus per nubem et locutus est ad eum auferens de spiritu qui erat in Mosen et dans septuaginta uiris cumque requieuisset in eis spiritus prophetauerunt nec ultra cessarunt 11:26 remanserant autem in castris duo uiri 11:26 But there remained two of the men in the camp, the

name of the one was Eldad, and the name of the other Medad: and the spirit rested upon them; and they were of them that were written, but went not out unto the tabernacle: and they prophesied in the camp.

quorum unus uocabatur Heldad et alter Medad super quos requieuit spiritus nam et ipsi descripti fuerant et non exierant ad tabernaculum 11:27 cumque prophetarent in ca- 11:27 And there ran a young man, and told Moses, and

said, Eldad and Medad do prophesy in the camp.

stris cucurrit puer et nuntiauit Mosi dicens Heldad et Medad prophetant in castris 11:28 statim Iosue filius Nun minis- 11:28 And Joshua the son of Nun, the servant of Moses,

one of his young men, answered and said, My lord Moses, forbid them.

ter Mosi et electus e pluribus ait domine mi Moses prohibe eos 11:29 at ille quid inquit aemularis pro me quis tribuat ut 11:29 And Moses said unto him, Enviest thou for my sake?

would God that all the LORD’s people were prophets, and that the LORD would put his spirit upon them!

Biblia Sacra Vulgata

omnis populus prophetet et det eis Dominus spiritum suum

11:30 And Moses gat him into the camp, he and the elders 11:30 reuersusque est Moses et maiores natu Israhel in castra

of Israel.

11:31 And there went forth a wind from the LORD, and 11:31 uentus autem egrediens a Domino arreptas trans mare

brought quails from the sea, and let them fall by the camp, as it were a day’s journey on this side, and as it were a day’s journey on the other side, round about the camp, and as it were two cubits high upon the face of the earth.

coturnices detulit et dimisit in castra itinere quantum uno die confici potest ex omni parte castrorum per circuitum uola-

11:32 And the people stood up all that day, and all that bantque in aere duobus cubitis altitudine super terram 11:32

night, and all the next day, and they gathered the quails: he that gathered least gathered ten homers: and they spread

them all abroad for themselves round about the camp. surgens ergo populus toto die illo et nocte ac die altero con-

gregauit coturnicum qui parum decem choros et siccauerunt

11:33 And while the flesh was yet between their teeth, ere eas per gyrum castrorum 11:33 adhuc carnes erant in deit was chewed, the wrath of the LORD was kindled against the people, and the LORD smote the people with a very

great plague. ntibus eorum nec defecerat huiuscemodi cibus et ecce furor

Domini concitatus in populum percussit eum plaga magna

11:34 And he called the name of that place Kibrothhat- nimis 11:34 uocatusque est ille locus sepulchra Concupis-

taavah: because there they buried the people that lusted. And the people journeyed from Kibrothhattaavah unto

Hazeroth; and abode at Hazeroth. centiae ibi enim sepelierunt populum qui desiderauerat egressi autem de sepulchris Concupiscentiae uenerunt in Aseroth et manserunt ibi

12:1 And Miriam and Aaron spake against Moses because 12:1 locutaque est Maria et Aaron contra Mosen propter

of the Ethiopian woman whom he had married: for he had married an Ethiopian woman.

uxorem eius aethiopissam 12:2 et dixerunt num per solum

12:2 And they said, Hath the LORD indeed spoken only by Moses? hath he not spoken also by us? And the LORD heard it.

Mosen locutus est Dominus nonne et nobis similiter est lo-

12:3 (Now the man Moses was very meek, above all the cutus quod cum audisset Dominus 12:3 erat enim Moses

men which were upon the face of the earth.)

uir mitissimus super omnes homines qui morabantur in terra

12:4 And the LORD spake suddenly unto Moses, and unto 12:4 statim locutus est ad eum et ad Aaron et Mariam eg-

Aaron, and unto Miriam, Come out ye three unto the tabernacle of the congregation. And they three came out.

redimini uos tantum tres ad tabernaculum foederis cumque

12:5 And the LORD came down in the pillar of the cloud, fuissent egressi 12:5 descendit Dominus in columna nubis

and stood in the door of the tabernacle, and called Aaron and Miriam: and they both came forth.

et stetit in introitu tabernaculi uocans Aaron et Mariam qui

12:6 And he said, Hear now my words: If there be a cum issent 12:6 dixit ad eos audite sermones meos si quis

prophet among you, I the LORD will make myself known unto him in a vision, and will speak unto him in a dream.

fuerit inter uos propheta Domini in uisione apparebo ei uel

12:7 My servant Moses is not so, who is faithful in all per somnium loquar ad illum 12:7 at non talis seruus meus

mine house. 12:8 With him will I speak mouth to mouth, even appar- Moses qui in omni domo mea fidelissimus est 12:8 ore ently, and not in dark speeches; and the similitude of the LORD shall he behold: wherefore then were ye not afraid

to speak against my servant Moses? enim ad os loquor ei et palam non per enigmata et figuras

Dominum uidet quare igitur non timuistis detrahere seruo

12:9 And the anger of the LORD was kindled against meo Mosi 12:9 iratusque contra eos abiit 12:10 nubes quo-

them; and he departed. 12:10 And the cloud departed from off the tabernacle; and, behold, Miriam became leprous, white as snow: and Aaron looked upon Miriam, and, behold, she was leprous.

que recessit quae erat super tabernaculum et ecce Maria apparuit candens lepra quasi nix cumque respexisset eam Aaron et uidisset perfusam lepra 12:11 ait ad Mosen obsecro 12:11 And Aaron said unto Moses, Alas, my lord, I be-

seech thee, lay not the sin upon us, wherein we have done foolishly, and wherein we have sinned.

domine mi ne inponas nobis hoc peccatum quod stulte commisimus 12:12 ne fiat haec quasi mortua et ut abortiuum 12:12 Let her not be as one dead, of whom the flesh is half

consumed when he cometh out of his mother’s womb.

quod proicitur de uulua matris suae ecce iam medium carnis eius deuoratum est lepra 12:13 clamauitque Moses ad Do- 12:13 And Moses cried unto the LORD, saying, Heal her

now, O God, I beseech thee.

minum dicens Deus obsecro sana eam 12:14 cui respondit 12:14 And the LORD said unto Moses, If her father had

but spit in her face, should she not be ashamed seven days? let her be shut out from the camp seven days, and after that

Dominus si pater eius spuisset in faciem illius nonne debue- let her be received in again. rat saltem septem dierum rubore suffundi separetur septem diebus extra castra et postea reuocabitur 12:15 exclusa est 12:15 And Miriam was shut out from the camp seven days:

and the people journeyed not till Miriam was brought in again.

itaque Maria extra castra septem diebus et populus non est motus de loco illo donec reuocata est Maria

13:1 profectus est de Aseroth fixis tentoriis in deserto Pha- 13:1 And afterward the people removed from Hazeroth,

and pitched in the wilderness of Paran.

ran 13:2 ibi locutus est Dominus ad Mosen dicens 13:3 13:2 And the LORD spake unto Moses, saying,

13:3 Send thou men, that they may search the land of Canaan, which I give unto the children of Israel: of every tribe of their fathers shall ye send a man, every one a ruler among them.

mitte uiros qui considerent terram Chanaan quam daturus sum filiis Israhel singulos de singulis tribubus ex principibus 13:4 fecit Moses quod Dominus imperarat de deserto Pha- 13:4 And Moses by the commandment of the LORD sent

them from the wilderness of Paran: all those men were heads of the children of Israel.

ran mittens principes uiros quorum ista sunt nomina 13:5

13:5 And these were their names: of the tribe of Reuben, Shammua the son of Zaccur.

de tribu Ruben Semmua filium Zecchur 13:6 de tribu Sy- 13:6 Of the tribe of Simeon, Shaphat the son of Hori. meon Saphat filium Huri 13:7 de tribu Iuda Chaleb filium 13:7 Of the tribe of Judah, Caleb the son of Jephunneh. Iepphonne 13:8 de tribu Isachar Igal filium Ioseph 13:9 13:8 Of the tribe of Issachar, Igal the son of Joseph.

13:9 Of the tribe of Ephraim, Oshea the son of Nun.

de tribu Ephraim Osee filium Nun 13:10 de tribu Beniamin _{13:10 Of the tribe of Benjamin, Palti the son of Raphu.} Phalti filium Raphu 13:11 de tribu Zabulon Geddihel filium 13:11 Of the tribe of Zebulun, Gaddiel the son of Sodi.

Sodi 13:12 de tribu Ioseph sceptri Manasse Gaddi filium 13:12 Of the tribe of Joseph, namely, of the tribe of Man-

asseh, Gaddi the son of Susi.

Susi 13:13 de tribu Dan Ammihel filium Gemalli 13:14 13:13 Of the tribe of Dan, Ammiel the son of Gemalli.

13:14 Of the tribe of Asher, Sethur the son of Michael.

de tribu Aser Sthur filium Michahel 13:15 de tribu Nept- _{13:15 Of the tribe of Naphtali, Nahbi the son of Vophsi.} hali Naabbi filium Vaphsi 13:16 de tribu Gad Guhel filium 13:16 Of the tribe of Gad, Geuel the son of Machi. Machi 13:17 haec sunt nomina uirorum quos misit Moses 13:17 These are the names of the men which Moses sent

to spy out the land. And Moses called Oshea the son of Nun Jehoshua.

ad considerandam terram uocauitque Osee filium Nun Iosue 13:18 misit ergo eos Moses ad considerandam terram Cha- 13:18 And Moses sent them to spy out the land of Canaan,

and said unto them, Get you up this way southward, and go up into the mountain:

naan et dixit ad eos ascendite per meridianam plagam cumque ueneritis ad montes 13:19 considerate terram qualis sit 13:19 And see the land, what it is, and the people that

dwelleth therein, whether they be strong or weak, few or many;

et populum qui habitator est eius utrum fortis sit an infirmus pauci numero an plures 13:20 ipsa terra bona an mala ur- 13:20 And what the land is that they dwell in, whether it

be good or bad; and what cities they be that they dwell in, whether in tents, or in strong holds;

Biblia Sacra Vulgata

13:21 And what the land is, whether it be fat or lean, bes quales muratae an absque muris 13:21 humus pinguis

whether there be wood therein, or not. And be ye of good courage, and bring of the fruit of the land. Now the time

was the time of the firstripe grapes. an sterilis nemorosa an absque arboribus confortamini et ad-

ferte nobis de fructibus terrae erat autem tempus quando iam

13:22 So they went up, and searched the land from the praecoquae uuae uesci possunt 13:22 cumque ascendissent

wilderness of Zin unto Rehob, as men come to Hamath.

explorauerunt terram a deserto Sin usque Roob intrantibus

13:23 And they ascended by the south, and came unto He- Emath 13:23 ascenderuntque ad meridiem et uenerunt in

bron; where Ahiman, Sheshai, and Talmai, the children of Anak, were. (Now Hebron was built seven years before

Zoan in Egypt.) Hebron ubi erant Ahiman et Sisai et Tholmai filii Enach nam

Hebron septem annis ante Tanim urbem Aegypti condita est

13:24 And they came unto the brook of Eshcol, and cut 13:24 pergentesque usque ad torrentem Botri absciderunt pal-

down from thence a branch with one cluster of grapes, and they bare it between two upon a staff; and they brought of

the pomegranates, and of the figs. mitem cum uua sua quem portauerunt in uecte duo uiri de 13:25 The place was called the brook Eshcol, because of malis quoque granatis et de ficis loci illius tulerunt 13:25

the cluster of grapes which the children of Israel cut down from thence.

qui appellatus est Neelescol id est torrens Botri eo quod bot-

13:26 And they returned from searching of the land after rum inde portassent filii Israhel 13:26 reuersique explorato-

forty days.

13:27 And they went and came to Moses, and to Aaron, res terrae post quadraginta dies omni regione circuita 13:27

and to all the congregation of the children of Israel, unto the wilderness of Paran, to Kadesh; and brought back word unto them, and unto all the congregation, and shewed them the fruit of the land.

uenerunt ad Mosen et Aaron et ad omnem coetum filiorum Israhel in desertum Pharan quod est in Cades locutique eis et

13:28 And they told him, and said, We came unto the land omni multitudini ostenderunt fructus terrae 13:28 et narra-

whither thou sentest us, and surely it floweth with milk and honey; and this is the fruit of it.

uerunt dicentes uenimus in terram ad quam misisti nos quae re uera fluit lacte et melle ut ex his fructibus cognosci potest

13:29 Nevertheless the people be strong that dwell in the 13:29 sed cultores fortissimos habet et urbes grandes atque

land, and the cities are walled, and very great: and moreover we saw the children of Anak there.

muratas stirpem Enach uidimus ibi 13:30 Amalech habitat

13:30 The Amalekites dwell in the land of the south: and the Hittites, and the Jebusites, and the Amorites, dwell in the mountains: and the Canaanites dwell by the sea, and by the coast of Jordan.

in meridie Hettheus et Iebuseus et Amorreus in montanis Chananeus uero moratur iuxta mare et circa fluenta Iorda-

13:31 And Caleb stilled the people before Moses, and said, nis 13:31 inter haec Chaleb conpescens murmur populi qui Let us go up at once, and possess it; for we are well able to overcome it.

oriebatur contra Mosen ait ascendamus et possideamus ter-

13:32 But the men that went up with him said, We be not ram quoniam poterimus obtinere eam 13:32 alii uero qui

able to go up against the people; for they are stronger than we.

fuerant cum eo dicebant nequaquam ad hunc populum uale-

13:33 And they brought up an evil report of the land which mus ascendere quia fortior nobis est 13:33 detraxeruntque

they had searched unto the children of Israel, saying, The land, through which we have gone to search it, is a land that eateth up the inhabitants thereof; and all the people that we saw in it are men of a great stature.

terrae quam inspexerant apud filios Israhel dicentes terram quam lustrauimus deuorat habitatores suos populum quem

13:34 And there we saw the giants, the sons of Anak, aspeximus procerae staturae est 13:34 ibi uidimus monstra

which come of the giants: and we were in our own sight as grasshoppers, and so we were in their sight.

quaedam filiorum Enach de genere giganteo quibus conparati quasi lucustae uidebamur 14:1 igitur uociferans omnis turba fleuit nocte illa 14:2 14:1 And all the congregation lifted up their voice, and

cried; and the people wept that night.

14:2 And all the children of Israel murmured against Moses and against Aaron: and the whole congregation said unto them, Would God that we had died in the land of Egypt! or would God we had died in this wilderness!

et murmurati sunt contra Mosen et Aaron cuncti filii Israhel dicentes 14:3 utinam mortui essemus in Aegypto et non in

14:3 And wherefore hath the LORD brought us unto this land, to fall by the sword, that our wives and our children should be a prey? were it not better for us to return into Egypt?

hac uasta solitudine utinam pereamus et non inducat nos Dominus in terram istam ne cadamus gladio et uxores ac liberi nostri ducantur captiui nonne melius est reuerti in Aegyptum 14:4 dixeruntque alter ad alterum constituamus nobis ducem 14:4 And they said one to another, Let us make a captain,

and let us return into Egypt.

et reuertamur in Aegyptum 14:5 quo audito Moses et Aaron 14:5 Then Moses and Aaron fell on their faces before all

the assembly of the congregation of the children of Israel.

ceciderunt proni in terram coram omni multitudine filiorum Israhel 14:6 at uero Iosue filius Nun et Chaleb filius Iep- 14:6 And Joshua the son of Nun, and Caleb the son of

Jephunneh, which were of them that searched the land, rent their clothes:

phonne qui et ipsi lustrauerant terram sciderunt uestimenta sua 14:7 et ad omnem multitudinem filiorum Israhel locuti 14:7 And they spake unto all the company of the children

of Israel, saying, The land, which we passed through to search it, is an exceeding good land.

sunt terram quam circuiuimus ualde bona est 14:8 si propi-

14:8 If the LORD delight in us, then he will bring us into this land, and give it us; a land which floweth with milk and honey.

tius fuerit Dominus inducet nos in eam et tradet humum lacte et melle manantem 14:9 nolite rebelles esse contra Domi- 14:9 Only rebel not ye against the LORD, neither fear ye

the people of the land; for they are bread for us: their defence is departed from them, and the LORD is with us:

num neque timeatis populum terrae huius quia sicut panem fear them not. ita eos possumus deuorare recessit ab illis omne praesidium Dominus nobiscum est nolite metuere 14:10 cumque cla- 14:10 But all the congregation bade stone them with

stones. And the glory of the LORD appeared in the tabernacle of the congregation before all the children of Israel.

maret omnis multitudo et lapidibus eos uellet opprimere apparuit gloria Domini super tectum foederis cunctis filiis Israhel 14:11 et dixit Dominus ad Mosen usquequo detrahet 14:11 And the LORD said unto Moses, How long will this

people provoke me? and how long will it be ere they believe me, for all the signs which I have shewed among

mihi populus iste quousque non credent mihi in omnibus si- them? gnis quae feci coram eis 14:12 feriam igitur eos pestilentia 14:12 I will smite them with the pestilence, and disinherit

them, and will make of thee a greater nation and mightier than they.

atque consumam te autem faciam principem super gentem magnam et fortiorem quam haec est 14:13 et ait Moses ad 14:13 And Moses said unto the LORD, Then the Egyp-

tians shall hear it, (for thou broughtest up this people in thy might from among them;)

Dominum ut audiant Aegyptii de quorum medio eduxisti populum istum 14:14 et habitatores terrae huius qui audierunt 14:14 And they will tell it to the inhabitants of this land:

for they have heard that thou LORD art among this people, that thou LORD art seen face to face, and that thy cloud standeth over them, and that thou goest before them, by day time in a pillar of a cloud, and in a pillar of fire by night.

quod tu Domine in populo isto sis et facie uidearis ad faciem et nubes tua protegat illos et in columna nubis praecedas eos per diem et in columna ignis per noctem 14:15 14:15 Now if thou shalt kill all this people as one man,

then the nations which have heard the fame of thee will speak, saying,

quod occideris tantam multitudinem quasi unum hominem et dicant 14:16 non poterat introducere populum in terram 14:16 Because the LORD was not able to bring this people

into the land which he sware unto them, therefore he hath slain them in the wilderness.

pro qua iurauerat idcirco occidit eos in solitudine 14:17 ma-

14:17 And now, I beseech thee, let the power of my LORD be great, according as thou hast spoken, saying,

Biblia Sacra Vulgata

14:18 The LORD is longsuffering, and of great mercy, for- gnificetur ergo fortitudo Domini sicut iurasti dicens 14:18

giving iniquity and transgression, and by no means clearing the guilty, visiting the iniquity of the fathers upon the

children unto the third and fourth generation. Dominus patiens et multae misericordiae auferens iniquitatem et scelera nullumque innoxium derelinquens qui uisitas peccata patrum in filios in tertiam et quartam generationem

14:19 Pardon, I beseech thee, the iniquity of this people 14:19 dimitte obsecro peccatum populi tui huius secundum

according unto the greatness of thy mercy, and as thou hast forgiven this people, from Egypt even until now.

magnitudinem misericordiae tuae sicut propitius fuisti egre-

14:20 And the LORD said, I have pardoned according to dientibus de Aegypto usque ad locum istum 14:20 dixitque

thy word: 14:21 But as truly as I live, all the earth shall be filled with Dominus dimisi iuxta uerbum tuum 14:21 uiuo ego et imthe glory of the LORD.

14:22 Because all those men which have seen my glory, plebitur gloria Domini uniuersa terra 14:22 attamen omnes

and my miracles, which I did in Egypt and in the wilderness, and have tempted me now these ten times, and have

not hearkened to my voice; homines qui uiderunt maiestatem meam et signa quae feci

in Aegypto et in solitudine et temptauerunt me iam per de-

14:23 Surely they shall not see the land which I sware unto cem uices nec oboedierunt uoci meae 14:23 non uidebunt their fathers, neither shall any of them that provoked me see it:

terram pro qua iuraui patribus eorum nec quisquam ex illis

14:24 But my servant Caleb, because he had another spirit qui detraxit mihi intuebitur eam 14:24 seruum meum Chawith him, and hath followed me fully, him will I bring into the land whereinto he went; and his seed shall possess it.

leb qui plenus alio spiritu secutus est me inducam in terram

14:25 (Now the Amalekites and the Canaanites dwelt in hanc quam circuiuit et semen eius possidebit eam 14:25

the valley.) Tomorrow turn you, and get you into the wilderness by the way of the Red sea.

quoniam Amalechites et Chananeus habitant in uallibus cras mouete castra et reuertimini in solitudinem per uiam maris

14:26 And the LORD spake unto Moses and unto Aaron, Rubri 14:26 locutusque est Dominus ad Mosen et Aaron

saying, 14:27 How long shall I bear with this evil congregation, dicens 14:27 usquequo multitudo haec pessima murmurat which murmur against me? I have heard the murmurings of the children of Israel, which they murmur against me.

contra me querellas filiorum Israhel audiui 14:28 dic ergo

14:28 Say unto them, As truly as I live, saith the LORD, as ye have spoken in mine ears, so will I do to you:

eis uiuo ego ait Dominus sicut locuti estis audiente me sic

14:29 Your carcases shall fall in this wilderness; and all faciam uobis 14:29 in solitudine hac iacebunt cadauera ues-

that were numbered of you, according to your whole number, from twenty years old and upward which have mur-

mured against me. tra omnes qui numerati estis a uiginti annis et supra et mur-14:30 Doubtless ye shall not come into the land, concern- murastis contra me 14:30 non intrabitis terram super quam

ing which I sware to make you dwell therein, save Caleb the son of Jephunneh, and Joshua the son of Nun.

leuaui manum meam ut habitare uos facerem praeter Chaleb

14:31 But your little ones, which ye said should be a prey, filium Iepphonne et Iosue filium Nun 14:31 paruulos autem them will I bring in, and they shall know the land which ye have despised.

uestros de quibus dixistis quod praedae hostibus forent intro-

14:32 But as for you, your carcases, they shall fall in this ducam ut uideant terram quae uobis displicuit 14:32 uestra

wilderness.

14:33 And your children shall wander in the wilderness cadauera iacebunt in solitudine 14:33 filii uestri erunt uagi

forty years, and bear your whoredoms, until your carcases be wasted in the wilderness.

in deserto annis quadraginta et portabunt fornicationem ue-

14:34 After the number of the days in which ye searched stram donec consumantur cadauera patrum in deserto 14:34

the land, even forty days, each day for a year, shall ye bear your iniquities, even forty years, and ye shall know my

breach of promise. iuxta numerum quadraginta dierum quibus considerastis terram annus pro die inputabitur et quadraginta annis recipietis iniquitates uestras et scietis ultionem meam 14:35 quoniam 14:35 I the LORD have said, I will surely do it unto all

this evil congregation, that are gathered together against me: in this wilderness they shall be consumed, and there

sicut locutus sum ita faciam omni multitudini huic pessi- they shall die. mae quae consurrexit aduersum me in solitudine hac deficiet et morietur 14:36 igitur omnes uiri quos miserat Mo- 14:36 And the men, which Moses sent to search the land,

who returned, and made all the congregation to murmur against him, by bringing up a slander upon the land,

ses ad contemplandam terram et qui reuersi murmurare fecerant contra eum omnem multitudinem detrahentes terrae quod esset mala 14:37 mortui sunt atque percussi in con- 14:37 Even those men that did bring up the evil report

upon the land, died by the plague before the LORD.

spectu Domini 14:38 Iosue autem filius Nun et Chaleb filius 14:38 But Joshua the son of Nun, and Caleb the son of

Jephunneh, which were of the men that went to search the land, lived still.

Iepphonne uixerunt ex omnibus qui perrexerant ad considerandam terram 14:39 locutusque est Moses uniuersa uerba 14:39 And Moses told these sayings unto all the children

of Israel: and the people mourned greatly.

haec ad omnes filios Israhel et luxit populus nimis 14:40 14:40 And they rose up early in the morning, and gat

them up into the top of the mountain, saying, Lo, we be here, and will go up unto the place which the LORD hath

et ecce mane primo surgentes ascenderunt uerticem montis promised: for we have sinned. atque dixerunt parati sumus ascendere ad locum de quo Dominus locutus est quia peccauimus 14:41 quibus Moses cur 14:41 And Moses said, Wherefore now do ye transgress

the commandment of the LORD? but it shall not prosper.

inquit transgredimini uerbum Domini quod uobis non cedet in prosperum 14:42 nolite ascendere non enim est Dominus 14:42 Go not up, for the LORD is not among you; that ye

be not smitten before your enemies.

uobiscum ne corruatis coram inimicis uestris 14:43 Ama- 14:43 For the Amalekites and the Canaanites are there be-

fore you, and ye shall fall by the sword: because ye are turned away from the LORD, therefore the LORD will not

lechites et Chananeus ante uos sunt quorum gladio corrue- be with you. tis eo quod nolueritis adquiescere Domino nec erit Dominus uobiscum 14:44 at illi contenebrati ascenderunt in uerticem 14:44 But they presumed to go up unto the hill top: never-

theless the ark of the covenant of the LORD, and Moses, departed not out of the camp.

montis arca autem testamenti Domini et Moses non recesserunt de castris 14:45 descenditque Amalechites et Chana- 14:45 Then the Amalekites came down, and the Canaan-

ites which dwelt in that hill, and smote them, and discomfited them, even unto Hormah.

neus qui habitabant in monte et percutiens eos atque concidens persecutus est usque Horma

15:1 locutus est Dominus ad Mosen dicens 15:2 loquere 15:1 And the LORD spake unto Moses, saying,

15:2 Speak unto the children of Israel, and say unto them, When ye be come into the land of your habitations, which I give unto you,

ad filios Israhel et dices ad eos cum ingressi fueritis terram habitationis uestrae quam ego dabo uobis 15:3 et feceritis 15:3 And will make an offering by fire unto the LORD, a

burnt offering, or a sacrifice in performing a vow, or in a freewill offering, or in your solemn feasts, to make a sweet

oblationem Domino in holocaustum aut uictimam uota sol- savour unto the LORD, of the herd or of the flock: uentes uel sponte offerentes munera aut in sollemnitatibus uestris adolentes odorem suauitatis Domino de bubus siue de ouibus 15:4 offeret quicumque immolauerit uictimam 15:4 Then shall he that offereth his offering unto the

LORD bring a meat offering of a tenth deal of flour mingled with the fourth part of an hin of oil.

sacrificium similae decimam partem oephi conspersae oleo

15:5 And the fourth part of an hin of wine for a drink offer- quod mensuram habebit quartam partem hin 15:5 et uinum ing shalt thou prepare with the burnt offering or sacrifice, for one lamb.

ad liba fundenda eiusdem mensurae dabit in holocausto siue

15:6 Or for a ram, thou shalt prepare for a meat offering in uictima per agnos singulos 15:6 et arietis erit sacrificium

two tenth deals of flour mingled with the third part of an hin of oil.

similae duarum decimarum quae conspersa sit oleo tertiae

15:7 And for a drink offering thou shalt offer the third part partis hin 15:7 et uinum ad libamentum tertiae partis eius-

of an hin of wine, for a sweet savour unto the LORD.

15:8 And when thou preparest a bullock for a burnt of- dem mensurae offeret in odorem suauitatis Domino 15:8

fering, or for a sacrifice in performing a vow, or peace offerings unto the LORD:

quando uero de bubus feceris holocaustum aut hostiam ut

15:9 Then shall he bring with a bullock a meat offering of impleas uotum uel pacificas uictimas 15:9 dabis per sin-

three tenth deals of flour mingled with half an hin of oil.

gulos boues similae tres decimas conspersae oleo quod ha-

15:10 And thou shalt bring for a drink offering half an hin beat medium mensurae hin 15:10 et uinum ad liba fundenda of wine, for an offering made by fire, of a sweet savour unto the LORD.

eiusdem mensurae in oblationem suauissimi odoris Domino

15:11 Thus shall it be done for one bullock, or for one ram, 15:11 sic facietis 15:12 per singulos boues et arietes et agnos

or for a lamb, or a kid.

15:12 According to the number that ye shall prepare, so shall ye do to every one according to their number.

et hedos 15:13 tam indigenae quam peregrini 15:14 eodem

15:13 All that are born of the country shall do these things after this manner, in offering an offering made by fire, of a sweet savour unto the LORD.

15:14 And if a stranger sojourn with you, or whosoever be among you in your generations, and will offer an offering made by fire, of a sweet savour unto the LORD; as ye do, so he shall do.

ritu offerent sacrificia 15:15 unum praeceptum erit atque

15:15 One ordinance shall be both for you of the congregation, and also for the stranger that sojourneth with you, an ordinance for ever in your generations: as ye are, so shall the stranger be before the LORD.

iudicium tam uobis quam aduenis terrae 15:16 locutus est

15:16 One law and one manner shall be for you, and for the stranger that sojourneth with you.

Dominus ad Mosen dicens 15:17 loquere filiis Israhel et di-

15:17 And the LORD spake unto Moses, saying,

ces ad eos 15:18 cum ueneritis in terram quam dabo uobis

15:18 Speak unto the children of Israel, and say unto them, When ye come into the land whither I bring you,

15:19 et comederitis de panibus regionis illius separabitis pri-

15:19 Then it shall be, that, when ye eat of the bread of the land, ye shall offer up an heave offering unto the LORD.

mitias Domino 15:20 de cibis uestris sicut de areis primitias

15:20 Ye shall offer up a cake of the first of your dough for an heave offering: as ye do the heave offering of the threshingfloor, so shall ye heave it.

separatis 15:21 ita et de pulmentis dabitis primitiua Domino

15:21 Of the first of your dough ye shall give unto the LORD an heave offering in your generations.

15:22 quod si per ignorantiam praeterieritis quicquam horum

15:22 And if ye have erred, and not observed all these commandments, which the LORD hath spoken unto Moses,

quae locutus est Dominus ad Mosen 15:23 et mandauit per

15:23 Even all that the LORD hath commanded you by the hand of Moses, from the day that the LORD commanded Moses, and henceforward among your generations;

eum ad uos a die qua coepit iubere et ultra 15:24 oblitaque

15:24 Then it shall be, if ought be committed by ignorance without the knowledge of the congregation, that all the congregation shall offer one young bullock for a burnt offering, for a sweet savour unto the LORD, with his meat offering, and his drink offering, according to the manner, and one kid of the goats for a sin offering.

fuerit facere multitudo offeret uitulum de armento holocaustum in odorem suauissimum Domino et sacrificium eius ac liba ut caerimoniae postulant hircumque pro peccato 15:25

15:25 And the priest shall make an atonement for all the congregation of the children of Israel, and it shall be forgiven them; for it is ignorance: and they shall bring their offering, a sacrifice made by fire unto the LORD, and their sin offering before the LORD, for their ignorance:

et rogabit sacerdos pro omni multitudine filiorum Israhel et dimittetur eis quoniam non sponte peccauerunt nihilominus offerentes incensum Domino pro se et pro peccato atque er-

15:26 And it shall be forgiven all the congregation of the rore suo 15:26 et dimittetur uniuersae plebi filiorum Isra-

children of Israel, and the stranger that sojourneth among them; seeing all the people were in ignorance.

hel et aduenis qui peregrinantur inter uos quoniam culpa est

15:27 And if any soul sin through ignorance, then he shall omnis populi per ignorantiam 15:27 quod si anima una nebring a she goat of the first year for a sin offering.

sciens peccauerit offeret capram anniculam pro peccato suo 15:28 et deprecabitur pro ea sacerdos quod inscia peccaue- 15:28 And the priest shall make an atonement for the soul

that sinneth ignorantly, when he sinneth by ignorance before the LORD, to make an atonement for him; and it shall

rit coram Domino inpetrabitque ei ueniam et dimittetur illi be forgiven him. 15:29 tam indigenis quam aduenis una lex erit omnium qui 15:29 Ye shall have one law for him that sinneth through

ignorance, both for him that is born among the children of Israel, and for the stranger that sojourneth among them.

peccauerint ignorantes 15:30 anima uero quae per super-

15:30 But the soul that doeth ought presumptuously, whether he be born in the land, or a stranger, the same reproacheth the LORD; and that soul shall be cut off from among his people.

biam aliquid commiserit siue ciuis sit ille siue peregrinus quoniam aduersum Dominum rebellis fuit peribit de populo suo 15:31 uerbum enim Domini contempsit et praeceptum 15:31 Because he hath despised the word of the LORD,

and hath broken his commandment, that soul shall utterly be cut off; his iniquity shall be upon him.

illius fecit irritum idcirco delebitur et portabit iniquitatem suam 15:32 factum est autem cum essent filii Israhel in soli- 15:32 And while the children of Israel were in the wilder-

ness, they found a man that gathered sticks upon the sabbath day.

tudine et inuenissent hominem colligentem ligna in die sabbati 15:33 obtulerunt eum Mosi et Aaron et uniuersae mul- 15:33 And they that found him gathering sticks brought

him unto Moses and Aaron, and unto all the congregation.

titudini 15:34 qui recluserunt eum in carcerem nescientes 15:34 And they put him in ward, because it was not de-

clared what should be done to him.

quid super eo facere deberent 15:35 dixitque Dominus ad 15:35 And the LORD said unto Moses, The man shall be

surely put to death: all the congregation shall stone him with stones without the camp.

Mosen morte moriatur homo iste obruat eum lapidibus omnis turba extra castra 15:36 cumque eduxissent eum foras 15:36 And all the congregation brought him without the

camp, and stoned him with stones, and he died; as the LORD commanded Moses.

obruerunt lapidibus et mortuus est sicut praeceperat Dominus 15:37 dixit quoque Dominus ad Mosen 15:38 loquere 15:37 And the LORD spake unto Moses, saying,

15:38 Speak unto the children of Israel, and bid them that they make them fringes in the borders of their garments throughout their generations, and that they put upon the fringe of the borders a ribband of blue:

filiis Israhel et dices ad eos ut faciant sibi fimbrias per angulos palliorum ponentes in eis uittas hyacinthinas 15:39 quas _{15:39 And it shall be unto you for a fringe, that ye may}

look upon it, and remember all the commandments of the LORD, and do them; and that ye seek not after your own heart and your own eyes, after which ye use to go a whoring:

cum uiderint recordentur omnium mandatorum Domini nec sequantur cogitationes suas et oculos per res uarias fornicantes 15:40 sed magis memores praeceptorum Domini faciant 15:40 That ye may remember, and do all my command-

ments, and be holy unto your God.

ea sintque sancti Deo suo 15:41 ego Dominus Deus uester 15:41 I am the LORD your God, which brought you out

of the land of Egypt, to be your God: I am the LORD your God.

qui eduxi uos de terra Aegypti ut essem uester Deus

16:1 ecce autem Core filius Isaar filii Caath filii Leui et 16:1 Now Korah, the son of Izhar, the son of Kohath, the

son of Levi, and Dathan and Abiram, the sons of Eliab, and On, the son of Peleth, sons of Reuben, took men:

Dathan atque Abiram filii Heliab Hon quoque filius Pheleth de filiis Ruben 16:2 surrexerunt contra Mosen aliique filio- 16:2 And they rose up before Moses, with certain of the

children of Israel, two hundred and fifty princes of the assembly, famous in the congregation, men of renown:

rum Israhel ducenti quinquaginta uiri proceres synagogae et

qui tempore concilii per nomina uocabantur 16:3 cumque 16:3 And they gathered themselves together against Moses and against Aaron, and said unto them, Ye take too much upon you, seeing all the congregation are holy, every one of them, and the LORD is among them: wherefore then lift ye up yourselves above the congregation of the LORD?

stetissent aduersum Mosen et Aaron dixerunt sufficiat uobis quia omnis multitudo sanctorum est et in ipsis est Dominus cur eleuamini super populum Domini 16:4 quod cum au- 16:4 And when Moses heard it, he fell upon his face: disset Moses cecidit pronus in faciem 16:5 locutusque ad 16:5 And he spake unto Korah and unto all his company,

saying, Even to morrow the LORD will shew who are his, and who is holy; and will cause him to come near unto him: even him whom he hath chosen will he cause to come near unto him.

Biblia Sacra Vulgata

Core et ad omnem multitudinem mane inquit notum faciet Dominus qui ad se pertineant et sanctos adplicabit sibi et

16:6 This do; Take you censers, Korah, and all his com- quos elegerit adpropinquabunt ei 16:6 hoc igitur facite tol-

pany;

lat unusquisque turibula sua tu Core et omne concilium tuum

16:7 And put fire therein, and put incense in them before 16:7 et hausto cras igne ponite desuper thymiama coram Do-

the LORD to morrow: and it shall be that the man whom the LORD doth choose, he shall be holy: ye take too much

upon you, ye sons of Levi. mino et quemcumque elegerit ipse erit sanctus multum eri-16:8 And Moses said unto Korah, Hear, I pray you, ye gimini filii Leui 16:8 dixitque rursum ad Core audite filii

sons of Levi: 16:9 Seemeth it but a small thing unto you, that the God of Leui 16:9 num parum uobis est quod separauit uos Deus Israel hath separated you from the congregation of Israel, to bring you near to himself to do the service of the tabernacle of the LORD, and to stand before the congregation to minister unto them?

Israhel ab omni populo et iunxit sibi ut seruiretis ei in cultu tabernaculi et staretis coram frequentia populi et ministrare-

16:10 And he hath brought thee near to him, and all thy tis ei 16:10 idcirco ad se fecit accedere te et omnes fratres

brethren the sons of Levi with thee: and seek ye the priesthood also?

tuos filios Leui ut uobis etiam sacerdotium uindicetis 16:11

16:11 For which cause both thou and all thy company are gathered together against the LORD: and what is Aaron, that ye murmur against him?

et omnis globus tuus stet contra Dominum quid est enim Aa-

16:12 And Moses sent to call Dathan and Abiram, the sons ron ut murmuretis contra eum 16:12 misit ergo Moses ut of Eliab: which said, We will not come up:

uocaret Dathan et Abiram filios Heliab qui responderunt non

16:13 Is it a small thing that thou hast brought us up out uenimus 16:13 numquid parum est tibi quod eduxisti nos of a land that floweth with milk and honey, to kill us in the wilderness, except thou make thyself altogether a prince

over us? de terra quae lacte et melle manabat ut occideres in deserto 16:14 Moreover thou hast not brought us into a land that nisi et dominatus fueris nostri 16:14 re uera induxisti nos in

floweth with milk and honey, or given us inheritance of fields and vineyards: wilt thou put out the eyes of these

men? we will not come up. terram quae fluit riuis lactis et mellis et dedisti nobis posses-

siones agrorum et uinearum an et oculos nostros uis eruere

16:15 And Moses was very wroth, and said unto the non uenimus 16:15 iratusque Moses ualde ait ad Dominum

LORD, Respect not thou their offering: I have not taken one ass from them, neither have I hurt one of them.

ne respicias sacrificia eorum tu scis quod ne asellum quidem umquam acceperim ab eis nec adflixerim quempiam eorum

16:16 And Moses said unto Korah, Be thou and all thy 16:16 dixitque ad Core tu et omnis congregatio tua state seor-

company before the LORD, thou, and they, and Aaron, to morrow:

sum coram Domino et Aaron die crastino separatim 16:17

16:17 And take every man his censer, and put incense in them, and bring ye before the LORD every man his censer, two hundred and fifty censers; thou also, and Aaron, each of you his censer.

tollite singuli turibula uestra et ponite super ea incensum offerentes Domino ducenta quinquaginta turibula Aaron quo-

16:18 And they took every man his censer, and put fire in que teneat turibulum suum 16:18 quod cum fecissent stan-

them, and laid incense thereon, and stood in the door of the tabernacle of the congregation with Moses and Aaron.

tibus Mosen et Aaron 16:19 et coaceruassent aduersum eos

16:19 And Korah gathered all the congregation against them unto the door of the tabernacle of the congregation: and the glory of the LORD appeared unto all the congregation.

omnem multitudinem ad ostium tabernaculi apparuit cunctis

16:20 And the LORD spake unto Moses and unto Aaron, gloria Domini 16:20 locutusque Dominus ad Mosen et Aa-

saying, 16:21 Separate yourselves from among this congregation, ron ait 16:21 separamini de medio congregationis huius ut that I may consume them in a moment.

16:22 And they fell upon their faces, and said, O God, the eos repente disperdam 16:22 qui ceciderunt proni in faciem

God of the spirits of all flesh, shall one man sin, and wilt thou be wroth with all the congregation?

atque dixerunt fortissime Deus spirituum uniuersae carnis num uno peccante contra omnes tua ira desaeuiet 16:23 et 16:23 And the LORD spake unto Moses, saying, ait Dominus ad Mosen 16:24 praecipe uniuerso populo ut 16:24 Speak unto the congregation, saying, Get you up

from about the tabernacle of Korah, Dathan, and Abiram.

separetur a tabernaculis Core et Dathan et Abiram 16:25 16:25 And Moses rose up and went unto Dathan and Abi-

ram; and the elders of Israel followed him.

surrexitque Moses et abiit ad Dathan et Abiram et sequentibus eum senioribus Israhel 16:26 dixit ad turbam recedite 16:26 And he spake unto the congregation, saying, Depart,

I pray you, from the tents of these wicked men, and touch nothing of their’s, lest ye be consumed in all their sins.

a tabernaculis hominum impiorum et nolite tangere quae ad eos pertinent ne inuoluamini in peccatis eorum 16:27 cum- 16:27 So they gat up from the tabernacle of Korah,

Dathan, and Abiram, on every side: and Dathan and Abiram came out, and stood in the door of their tents, and their

que recessissent a tentoriis eorum per circuitum Dathan et wives, and their sons, and their little children. Abiram egressi stabant in introitu papilionum suorum cum uxoribus et liberis omnique frequentia 16:28 et ait Moses 16:28 And Moses said, Hereby ye shall know that the

LORD hath sent me to do all these works; for I have not done them of mine own mind.

in hoc scietis quod Dominus miserit me ut facerem uniuersa quae cernitis et non ex proprio ea corde protulerim 16:29 si 16:29 If these men die the common death of all men, or

if they be visited after the visitation of all men; then the LORD hath not sent me.

consueta hominum morte interierint et uisitauerit eos plaga qua et ceteri uisitari solent non misit me Dominus 16:30 16:30 But if the LORD make a new thing, and the earth

open her mouth, and swallow them up, with all that appertain unto them, and they go down quick into the pit; then ye shall understand that these men have provoked the LORD.

sin autem nouam rem fecerit Dominus ut aperiens terra os suum degluttiat eos et omnia quae ad illos pertinent descenderintque uiuentes in infernum scietis quod blasphemauerint Dominum 16:31 confestim igitur ut cessauit loqui dis- 16:31 And it came to pass, as he had made an end of speak-

ing all these words, that the ground clave asunder that was under them:

rupta est terra sub pedibus eorum 16:32 et aperiens os suum

16:32 And the earth opened her mouth, and swallowed them up, and their houses, and all the men that appertained unto Korah, and all their goods.

deuorauit illos cum tabernaculis suis et uniuersa substantia 16:33 descenderuntque uiui in infernum operti humo et perie- 16:33 They, and all that appertained to them, went down

alive into the pit, and the earth closed upon them: and they perished from among the congregation.

runt de medio multitudinis 16:34 at uero omnis Israhel qui

16:34 And all Israel that were round about them fled at the cry of them: for they said, Lest the earth swallow us up also.

stabat per gyrum fugit ad clamorem pereuntium dicens ne forte et nos terra degluttiat 16:35 sed et ignis egressus a Do- 16:35 And there came out a fire from the LORD, and con-

sumed the two hundred and fifty men that offered incense.

mino interfecit ducentos quinquaginta uiros qui offerebant incensum 16:36 locutusque est Dominus ad Mosen dicens 16:36 And the LORD spake unto Moses, saying, 16:37 praecipe Eleazaro filio Aaron sacerdotis ut tollat tu- 16:37 Speak unto Eleazar the son of Aaron the priest, that

he take up the censers out of the burning, and scatter thou the fire yonder; for they are hallowed.

ribula quae iacent in incendio et ignem huc illucque dispergat quoniam sanctificata sunt 16:38 in mortibus peccatorum 16:38 The censers of these sinners against their own souls,

let them make them broad plates for a covering of the altar: for they offered them before the LORD, therefore they are hallowed: and they shall be a sign unto the children of Israel.

producatque ea in lamminas et adfigat altari eo quod oblatum sit in eis incensum Domino et sanctificata sint ut cernant ea pro signo et monumento filii Israhel 16:39 tulit ergo Eleazar 16:39 And Eleazar the priest took the brasen censers,

wherewith they that were burnt had offered; and they were made broad plates for a covering of the altar:

Biblia Sacra Vulgata

sacerdos turibula aenea in quibus obtulerant hii quos incendium deuorauit et produxit ea in lamminas adfigens altari

16:40 To be a memorial unto the children of Israel, that no 16:40 ut haberent postea filii Israhel quibus commonerentur

stranger, which is not of the seed of Aaron, come near to offer incense before the LORD; that he be not as Korah, and as his company: as the LORD said to him by the hand of Moses.

ne quis accedat alienigena et qui non est de semine Aaron ad offerendum incensum Domino ne patiatur sicut passus est Core et omnis congregatio eius loquente Domino ad Mosen

16:41 But on the morrow all the congregation of the chil- 16:41 murmurauit autem omnis multitudo filiorum Israhel se-

dren of Israel murmured against Moses and against Aaron, saying, Ye have killed the people of the LORD.

quenti die contra Mosen et Aaron dicens uos interfecistis

16:42 And it came to pass, when the congregation was populum Domini 16:42 cumque oreretur seditio et tumul-

gathered against Moses and against Aaron, that they looked toward the tabernacle of the congregation: and, behold, the cloud covered it, and the glory of the LORD appeared.

tus incresceret 16:43 Moses et Aaron fugerunt ad taberna-

16:43 And Moses and Aaron came before the tabernacle of the congregation.

culum foederis quod postquam ingressi sunt operuit nubes

16:44 And the LORD spake unto Moses, saying, et apparuit gloria Domini 16:44 dixitque Dominus ad Mo-16:45 Get you up from among this congregation, that I sen 16:45 recedite de medio huius multitudinis etiam nunc

may consume them as in a moment. And they fell upon their faces.

delebo eos cumque iacerent in terra 16:46 dixit Moses ad

16:46 And Moses said unto Aaron, Take a censer, and put fire therein from off the altar, and put on incense, and go quickly unto the congregation, and make an atonement for them: for there is wrath gone out from the LORD; the plague is begun.

Aaron tolle turibulum et hausto igne de altari mitte incensum desuper pergens cito ad populum ut roges pro eis iam enim

16:47 And Aaron took as Moses commanded, and ran into egressa est ira a Domino et plaga desaeuit 16:47 quod cum

the midst of the congregation; and, behold, the plague was begun among the people: and he put on incense, and made

an atonement for the people. fecisset Aaron et cucurrisset ad mediam multitudinem quam 16:48 And he stood between the dead and the living; and iam uastabat incendium obtulit thymiama 16:48 et stans in-

the plague was stayed.

ter mortuos ac uiuentes pro populo deprecatus est et plaga

16:49 Now they that died in the plague were fourteen thou- cessauit 16:49 fuerunt autem qui percussi sunt quattuorde-

sand and seven hundred, beside them that died about the matter of Korah.

cim milia hominum et septingenti absque his qui perierant

16:50 And Aaron returned unto Moses unto the door of the in seditione Core 16:50 reuersusque est Aaron ad Mosen tabernacle of the congregation: and the plague was stayed.

ad ostium tabernaculi foederis postquam quieuit interitus

17:1 And the LORD spake unto Moses, saying, 17:1 et locutus est Dominus ad Mosen dicens 17:2 lo-

17:2 Speak unto the children of Israel, and take of every one of them a rod according to the house of their fathers, of all their princes according to the house of their fathers twelve rods: write thou every man’s name upon his rod.

quere ad filios Israhel et accipe ab eis uirgas singulas per cognationes suas a cunctis principibus tribuum uirgas duode-

17:3 And thou shalt write Aaron’s name upon the rod of cim et uniuscuiusque nomen superscribes uirgae suae 17:3

Levi: for one rod shall be for the head of the house of their fathers.

nomen autem Aaron erit in tribu Leui et una uirga cunctas

17:4 And thou shalt lay them up in the tabernacle of the eorum familias continebit 17:4 ponesque eas in tabernaculo

congregation before the testimony, where I will meet with you.

foederis coram testimonio ubi loquar ad te 17:5 quem ex his

17:5 And it shall come to pass, that the man’s rod, whom I shall choose, shall blossom: and I will make to cease from me the murmurings of the children of Israel, whereby they murmur against you.

elegero germinabit uirga eius et cohibebo a me querimonias

17:6 And Moses spake unto the children of Israel, and ev- filiorum Israhel quibus contra uos murmurant 17:6 locutus-

ery one of their princes gave him a rod apiece, for each prince one, according to their fathers’ houses, even twelve rods: and the rod of Aaron was among their rods.

que est Moses ad filios Israhel et dederunt ei omnes principes uirgas per singulas tribus fueruntque uirgae duodecim absque uirga Aaron 17:7 quas cum posuisset Moses coram Do- 17:7 And Moses laid up the rods before the LORD in the

tabernacle of witness.

mino in tabernaculo testimonii 17:8 sequenti die regressus 17:8 And it came to pass, that on the morrow Moses went

into the tabernacle of witness; and, behold, the rod of Aaron for the house of Levi was budded, and brought forth

inuenit germinasse uirgam Aaron in domo Leui et turgen- buds, and bloomed blossoms, and yielded almonds. tibus gemmis eruperant flores qui foliis dilatatis in amigdalas deformati sunt 17:9 protulit ergo Moses omnes uirgas 17:9 And Moses brought out all the rods from before the

LORD unto all the children of Israel: and they looked, and took every man his rod.

de conspectu Domini ad cunctos filios Israhel uideruntque et receperunt singuli uirgas suas 17:10 dixitque Dominus 17:10 And the LORD said unto Moses, Bring Aaron’s rod

again before the testimony, to be kept for a token against the rebels; and thou shalt quite take away their murmur-

ad Mosen refer uirgam Aaron in tabernaculum testimonii ut ings from me, that they die not. seruetur ibi in signum rebellium filiorum et quiescant querellae eorum a me ne moriantur 17:11 fecitque Moses sicut 17:11 And Moses did so: as the LORD commanded him,

so did he.

praeceperat Dominus 17:12 dixerunt autem filii Israhel ad 17:12 And the children of Israel spake unto Moses, saying,

Behold, we die, we perish, we all perish.

Mosen ecce consumpti sumus omnes periuimus 17:13 qui- 17:13 Whosoever cometh any thing near unto the taber-

nacle of the LORD shall die: shall we be consumed with dying?

cumque accedit ad tabernaculum Domini moritur num usque ad internicionem cuncti delendi sumus

18:1 dixitque Dominus ad Aaron tu et filii tui et domus 18:1 And the LORD said unto Aaron, Thou and thy sons

and thy father’s house with thee shall bear the iniquity of the sanctuary: and thou and thy sons with thee shall bear

patris tui tecum portabitis iniquitatem sanctuarii et tu et fi- the iniquity of your priesthood. lii tui simul sustinebitis peccata sacerdotii uestri 18:2 sed 18:2 And thy brethren also of the tribe of Levi, the tribe

of thy father, bring thou with thee, that they may be joined unto thee, and minister unto thee: but thou and thy sons

et fratres tuos de tribu Leui et sceptro patris tui sume te- with thee shall minister before the tabernacle of witness.

cum praestoque sint et ministrent tibi tu autem et filii tui ministrabitis in tabernaculo testimonii 18:3 excubabuntque 18:3 And they shall keep thy charge, and the charge of all

the tabernacle: only they shall not come nigh the vessels of the sanctuary and the altar, that neither they, nor ye also,

Leuitae ad praecepta tua et ad cuncta opera tabernaculi ita die. dumtaxat ut ad uasa sanctuarii et altare non accedant ne et illi moriantur et uos pereatis simul 18:4 sint autem tecum et 18:4 And they shall be joined unto thee, and keep the

charge of the tabernacle of the congregation, for all the service of the tabernacle: and a stranger shall not come

excubent in custodiis tabernaculi et in omnibus caerimoniis nigh unto you. eius alienigena non miscebitur uobis 18:5 excubate in cus- 18:5 And ye shall keep the charge of the sanctuary, and the

charge of the altar: that there be no wrath any more upon the children of Israel.

todia sanctuarii et in ministerio altaris ne oriatur indignatio super filios Israhel 18:6 ego dedi uobis fratres uestros Le- 18:6 And I, behold, I have taken your brethren the Levites

from among the children of Israel: to you they are given as a gift for the LORD, to do the service of the tabernacle

uitas de medio filiorum Israhel et tradidi donum Domino ut of the congregation. seruiant in ministeriis tabernaculi eius 18:7 tu autem et filii 18:7 Therefore thou and thy sons with thee shall keep your

priest’s office for everything of the altar, and within the vail; and ye shall serve: I have given your priest’s office unto you as a service of gift: and the stranger that cometh nigh shall be put to death.

tui custodite sacerdotium uestrum et omnia quae ad cultum altaris pertinent et intra uelum sunt per sacerdotes admini-

Biblia Sacra Vulgata

18:8 And the LORD spake unto Aaron, Behold, I also have strabuntur si quis externus accesserit occidetur 18:8 locutus

given thee the charge of mine heave offerings of all the hallowed things of the children of Israel; unto thee have I given them by reason of the anointing, and to thy sons, by an ordinance for ever.

est Dominus ad Aaron ecce dedi tibi custodiam primitiarum mearum omnia quae sanctificantur a filiis Israhel tibi tradidi

18:9 This shall be thine of the most holy things, reserved et filiis tuis pro officio sacerdotali legitima sempiterna 18:9

from the fire: every oblation of theirs, every meat offering of theirs, and every sin offering of theirs, and every trespass offering of theirs which they shall render unto me, shall be most holy for thee and for thy sons.

haec ergo accipies de his quae sanctificantur et oblata sunt Domino omnis oblatio et sacrificium et quicquid pro peccato atque delicto redditur mihi et cedet in sancta sancto-

18:10 In the most holy place shalt thou eat it; every male rum tuum erit et filiorum tuorum 18:10 in sanctuario come-

shall eat it: it shall be holy unto thee.

des illud mares tantum edent ex eo quia consecratum est tibi

18:11 And this is thine; the heave offering of their gift, 18:11 primitias autem quas uouerint et obtulerint filii Israhel

with all the wave offerings of the children of Israel: I have given them unto thee, and to thy sons and to thy daughters with thee, by a statute for ever: every one that is clean in thy house shall eat of it.

tibi dedi et filiis ac filiabus tuis iure perpetuo qui mundus

18:12 All the best of the oil, and all the best of the wine, est in domo tua uescetur eis 18:12 omnem medullam olei and of the wheat, the firstfruits of them which they shall offer unto the LORD, them have I given thee.

et uini ac frumenti quicquid offerunt primitiarum Domino

18:13 And whatsoever is first ripe in the land, which they tibi dedi 18:13 uniuersa frugum initia quas gignit humus et

shall bring unto the LORD, shall be thine; every one that is clean in thine house shall eat of it.

Domino deportantur cedent in usus tuos qui mundus est in

18:14 Every thing devoted in Israel shall be thine. domo tua uescetur eis 18:14 omne quod ex uoto reddide-18:15 Every thing that openeth the matrix in all flesh, rint filii Israhel tuum erit 18:15 quicquid primum erumpet

which they bring unto the LORD, whether it be of men or beasts, shall be thine: nevertheless the firstborn of man shalt thou surely redeem, and the firstling of unclean beasts shalt thou redeem.

e uulua cunctae carnis quam offerunt Domino siue ex hominibus siue de pecoribus fuerit tui iuris erit ita dumtaxat ut pro hominis primogenito pretium accipias et omne animal

18:16 And those that are to be redeemed from a month old quod inmundum est rediii facias 18:16 cuius redeeptio erit

shalt thou redeem, according to thine estimation, for the money of five shekels, after the shekel of the sanctuary,

which is twenty gerahs. post unum mensem siclis argenti quinque pondere sanctuarii 18:17 But the firstling of a cow, or the firstling of a sheep, siclus uiginti obolos habet 18:17 primogenitum autem bo-

or the firstling of a goat, thou shalt not redeem; they are holy: thou shalt sprinkle their blood upon the altar, and shalt burn their fat for an offering made by fire, for a sweet savour unto the LORD.

uis et ouis et caprae non facies rediii quia sanctificata sunt Domino sanguinem tantum eorum fundes super altare et ad-

18:18 And the flesh of them shall be thine, as the wave ipes adolebis in suauissimum odorem Domino 18:18 carnes

breast and as the right shoulder are thine.

uero in usum tuum cedent sicut pectusculum consecratum

18:19 All the heave offerings of the holy things, which the et armus dexter tua erunt 18:19 omnes primitias sanctuarii

children of Israel offer unto the LORD, have I given thee, and thy sons and thy daughters with thee, by a statute for ever: it is a covenant of salt for ever before the LORD unto thee and to thy seed with thee.

quas offerunt filii Israhel Domino tibi dedi et filiis ac filiabus tuis iure perpetuo pactum salis est sempiternum coram

18:20 And the LORD spake unto Aaron, Thou shalt have Domino tibi ac filiis tuis 18:20 dixitque Dominus ad Aaron

no inheritance in their land, neither shalt thou have any part among them: I am thy part and thine inheritance

among the children of Israel. in terra eorum nihil possidebitis nec habebitis partem inter 18:21 And, behold, I have given the children of Levi all the eos ego pars et hereditas tua in medio filiorum Israhel 18:21

tenth in Israel for an inheritance, for their service which they serve, even the service of the tabernacle of the congregation.

filiis autem Leui dedi omnes decimas Israhelis in possessionem pro ministerio quo seruiunt mihi in tabernaculo foederis 18:22 ut non accedant ultra filii Israhel ad tabernacu- 18:22 Neither must the children of Israel henceforth come

nigh the tabernacle of the congregation, lest they bear sin, and die.

lum nec committant peccatum mortiferum 18:23 solis filiis

18:23 But the Levites shall do the service of the tabernacle of the congregation, and they shall bear their iniquity: it shall be a statute for ever throughout your generations, that among the children of Israel they have no inheritance.

Leui mihi in tabernaculo seruientibus et portantibus peccata populi legitimum sempiternum erit in generationibus uestris nihil aliud possidebunt 18:24 decimarum oblatione contenti 18:24 But the tithes of the children of Israel, which they

offer as an heave offering unto the LORD, I have given to the Levites to inherit: therefore I have said unto them, Among the children of Israel they shall have no inheritance.

quas in usus eorum et necessaria separaui 18:25 locutusque

18:25 And the LORD spake unto Moses, saying,

est Dominus ad Mosen dicens 18:26 praecipe Leuitis atque

18:26 Thus speak unto the Levites, and say unto them, When ye take of the children of Israel the tithes which I have given you from them for your inheritance, then ye shall offer up an heave offering of it for the LORD, even a tenth part of the tithe.

denuntia cum acceperitis a filiis Israhel decimas quas dedi uobis primitias earum offerte Domino id est decimam partem decimae 18:27 ut reputetur uobis in oblationem primi- 18:27 And this your heave offering shall be reckoned unto

you, as though it were the corn of the threshingfloor, and as the fulness of the winepress.

tiuorum tam de areis quam de torcularibus 18:28 et uniuer-

18:28 Thus ye also shall offer an heave offering unto the LORD of all your tithes, which ye receive of the children of Israel; and ye shall give thereof the LORD’s heave offering to Aaron the priest.

sis quorum accipitis primitias offerte Domino et date Aaron sacerdoti 18:29 omnia quae offertis ex decimis et in donaria 18:29 Out of all your gifts ye shall offer every heave offer-

ing of the LORD, of all the best thereof, even the hallowed part thereof out of it.

Domini separatis optima et electa erunt 18:30 dicesque ad

18:30 Therefore thou shalt say unto them, When ye have heaved the best thereof from it, then it shall be counted unto the Levites as the increase of the threshingfloor, and as the increase of the winepress.

eos si praeclara et meliora quaeque obtuleritis ex decimis reputabitur uobis quasi de area et torculari dederitis primitias 18:31 et comedetis eas in omnibus locis uestris tam uos quam 18:31 And ye shall eat it in every place, ye and your house-

holds: for it is your reward for your service in the tabernacle of the congregation.

familiae uestrae quia pretium est pro ministerio quo seruitis in tabernaculo testimonii 18:32 et non peccabitis super hoc 18:32 And ye shall bear no sin by reason of it, when ye

have heaved from it the best of it: neither shall ye pollute the holy things of the children of Israel, lest ye die.

egregia uobis et pinguia reseruantes ne polluatis oblationes filiorum Israhel et moriamini

19:1 locutusque est Dominus ad Mosen et Aaron dicens 19:1 And the LORD spake unto Moses and unto Aaron,

saying,

19:2 ista est religio uictimae quam constituit Dominus prae- 19:2 This is the ordinance of the law which the LORD hath

commanded, saying, Speak unto the children of Israel, that they bring thee a red heifer without spot, wherein is no

cipe filiis Israhel ut adducant ad te uaccam rufam aetatis in- blemish, and upon which never came yoke:

tegrae in qua nulla sit macula nec portauerit iugum 19:3 19:3 And ye shall give her unto Eleazar the priest, that he may bring her forth without the camp, and one shall slay her before his face:

tradetisque eam Eleazaro sacerdoti qui eductam extra castra immolabit in conspectu omnium 19:4 et tinguens digi- 19:4 And Eleazar the priest shall take of her blood with

his finger, and sprinkle of her blood directly before the tabernacle of the congregation seven times:

tum in sanguine eius asperget contra fores tabernaculi septem uicibus 19:5 conburetque eam cunctis uidentibus tam 19:5 And one shall burn the heifer in his sight; her skin,

and her flesh, and her blood, with her dung, shall he burn:

pelle et carnibus eius quam sanguine et fimo flammae traditis 19:6 lignum quoque cedrinum et hysopum coccum- 19:6 And the priest shall take cedar wood, and hyssop,

and scarlet, and cast it into the midst of the burning of the heifer.

Biblia Sacra Vulgata

que bis tinctum sacerdos mittet in flammam quae uaccam

19:7 Then the priest shall wash his clothes, and he shall uorat 19:7 et tunc demum lotis uestibus et corpore suo in-

bathe his flesh in water, and afterward he shall come into the camp, and the priest shall be unclean until the even.

gredietur in castra commaculatusque erit usque ad uespe-

19:8 And he that burneth her shall wash his clothes in wa- ram 19:8 sed et ille qui conbuserit eam lauabit uestimenta

ter, and bathe his flesh in water, and shall be unclean until the even.

sua et corpus et inmundus erit usque ad uesperam 19:9 col-

19:9 And a man that is clean shall gather up the ashes of the heifer, and lay them up without the camp in a clean place, and it shall be kept for the congregation of the children of Israel for a water of separation: it is a purification for sin.

liget autem uir mundus cineres uaccae et effundet eos extra castra in loco purissimo ut sint multitudini filiorum Israhel in custodiam et in aquam aspersionis quia pro peccato

19:10 And he that gathereth the ashes of the heifer shall uacca conbusta est 19:10 cumque lauerit qui uaccae porta-

wash his clothes, and be unclean until the even: and it shall be unto the children of Israel, and unto the stranger

that sojourneth among them, for a statute for ever. uerat cineres uestimenta sua inmundus erit usque ad uespe-

rum habebunt hoc filii Israhel et aduenae qui habitant inter

19:11 He that toucheth the dead body of any man shall be eos sanctum iure perpetuo 19:11 qui tetigerit cadauer ho-

unclean seven days. 19:12 He shall purify himself with it on the third day, and minis et propter hoc septem diebus fuerit inmundus 19:12 on the seventh day he shall be clean: but if he purify not himself the third day, then the seventh day he shall not be

clean. aspergetur ex hac aqua die tertio et septimo et sic munda-

bitur si die tertio aspersus non fuerit septimo non poterit

19:13 Whosoever toucheth the dead body of any man that emundari 19:13 omnis qui tetigerit humanae animae mor-

is dead, and purifieth not himself, defileth the tabernacle of the LORD; and that soul shall be cut off from Israel: because the water of separation was not sprinkled upon him, he shall be unclean; his uncleanness is yet upon him.

ticinum et aspersus hac commixtione non fuerit polluet tabernaculum Domini et peribit ex Israhel quia aqua expiationis non est aspersus inmundus erit et manebit spurcitia eius

19:14 This is the law, when a man dieth in a tent: all that super eum 19:14 ista est lex hominis qui moritur in taberna-

come into the tent, and all that is in the tent, shall be unclean seven days.

culo omnes qui ingrediuntur tentorium illius et uniuersa uasa

19:15 And every open vessel, which hath no covering quae ibi sunt polluta erunt septem diebus 19:15 uas quod

bound upon it, is unclean.

non habuerit operculum nec ligaturam desuper inmundum

19:16 And whosoever toucheth one that is slain with a erit 19:16 si quis in agro tetigerit cadauer occisi hominis

sword in the open fields, or a dead body, or a bone of a man, or a grave, shall be unclean seven days.

aut per se mortui siue os illius uel sepulchrum inmundus erit

19:17 And for an unclean person they shall take of the septem diebus 19:17 tollent de cineribus conbustionis atque

ashes of the burnt heifer of purification for sin, and running water shall be put thereto in a vessel:

peccati et mittent aquas uiuas super eos in uas 19:18 in qui-

19:18 And a clean person shall take hyssop, and dip it in the water, and sprinkle it upon the tent, and upon all the vessels, and upon the persons that were there, and upon him that touched a bone, or one slain, or one dead, or a grave:

bus cum homo mundus tinxerit hysopum asperget eo omne tentorium et cunctam supellectilem et homines huiuscemodi

19:19 And the clean person shall sprinkle upon the un- contagione pollutos 19:19 atque hoc modo mundus lustra-

clean on the third day, and on the seventh day: and on the seventh day he shall purify himself, and wash his clothes,

and bathe himself in water, and shall be clean at even. bit inmundum tertio et septimo die expiatusque die septimo

lauabit et se et uestimenta sua et mundus erit ad uesperam

19:20 But the man that shall be unclean, and shall not 19:20 si quis hoc ritu non fuerit expiatus peribit anima il-

purify himself, that soul shall be cut off from among the congregation, because he hath defiled the sanctuary of the LORD: the water of separation hath not been sprinkled upon him; he is unclean.

lius de medio ecclesiae quia sanctuarium Domini polluit et non est aqua lustrationis aspersus 19:21 erit hoc praecep- 19:21 And it shall be a perpetual statute unto them, that

he that sprinkleth the water of separation shall wash his clothes; and he that toucheth the water of separation shall

tum legitimum sempiternum ipse quoque qui aspergit aquas be unclean until even. lauabit uestimenta sua omnis qui tetigerit aquas expiationis inmundus erit usque ad uesperam 19:22 quicquid tetigerit 19:22 And whatsoever the unclean person toucheth shall

be unclean; and the soul that toucheth it shall be unclean until even.

inmundus inmundum faciet et anima quae horum quippiam tetigerit inmunda erit usque ad uesperum

20:1 ueneruntque filii Israhel et omnis multitudo in deser- 20:1 Then came the children of Israel, even the whole con-

gregation, into the desert of Zin in the first month: and the people abode in Kadesh; and Miriam died there, and was

tum Sin mense primo et mansit populus in Cades mortua- buried there. que est ibi Maria et sepulta in eodem loco 20:2 cumque 20:2 And there was no water for the congregation: and

they gathered themselves together against Moses and against Aaron.

indigeret aqua populus coierunt aduersum Mosen et Aaron 20:3 et uersi in seditionem dixerunt utinam perissemus inter 20:3 And the people chode with Moses, and spake, say-

ing, Would God that we had died when our brethren died before the LORD!

fratres nostros coram Domino 20:4 cur eduxistis ecclesiam

20:4 And why have ye brought up the congregation of the LORD into this wilderness, that we and our cattle should die there?

Domini in solitudinem ut et nos et nostra iumenta moriantur 20:5 quare nos fecistis ascendere de Aegypto et adduxi- 20:5 And wherefore have ye made us to come up out of

Egypt, to bring us in unto this evil place? it is no place of seed, or of figs, or of vines, or of pomegranates; neither is

stis in locum istum pessimum qui seri non potest qui nec there any water to drink. ficum gignit nec uineas nec mala granata insuper et aquam non habet ad bibendum 20:6 ingressusque Moses et Aaron 20:6 And Moses and Aaron went from the presence of the

assembly unto the door of the tabernacle of the congregation, and they fell upon their faces: and the glory of the

dimissa multitudine tabernaculum foederis corruerunt proni LORD appeared unto them. in terram et apparuit gloria Domini super eos 20:7 locu- 20:7 And the LORD spake unto Moses, saying,

tusque est Dominus ad Mosen dicens 20:8 tolle uirgam et 20:8 Take the rod, and gather thou the assembly together, thou, and Aaron thy brother, and speak ye unto the rock before their eyes; and it shall give forth his water, and thou shalt bring forth to them water out of the rock: so thou shalt give the congregation and their beasts drink.

congrega populum tu et Aaron frater tuus et loquimini ad petram coram eis et illa dabit aquas cumque eduxeris aquam de petra bibet omnis multitudo et iumenta eius 20:9 tulit igitur 20:9 And Moses took the rod from before the LORD, as

he commanded him.

Moses uirgam quae erat in conspectu Domini sicut praeceperat ei 20:10 congregata multitudine ante petram dixitque 20:10 And Moses and Aaron gathered the congregation

together before the rock, and he said unto them, Hear now, ye rebels; must we fetch you water out of this rock?

eis audite rebelles et increduli num de petra hac uobis aquam poterimus eicere 20:11 cumque eleuasset Moses manum 20:11 And Moses lifted up his hand, and with his rod he

smote the rock twice: and the water came out abundantly, and the congregation drank, and their beasts also.

percutiens uirga bis silicem egressae sunt aquae largissimae ita ut et populus biberet et iumenta 20:12 dixitque Dominus 20:12 And the LORD spake unto Moses and Aaron, Be-

cause ye believed me not, to sanctify me in the eyes of the children of Israel, therefore ye shall not bring this congre-

ad Mosen et Aaron quia non credidistis mihi ut sanctificare- gation into the land which I have given them. tis me coram filiis Israhel non introducetis hos populos in terram quam dabo eis 20:13 haec est aqua Contradictionis 20:13 This is the water of Meribah; because the children

of Israel strove with the LORD, and he was sanctified in them.

Biblia Sacra Vulgata

ubi iurgati sunt filii Israhel contra Dominum et sanctifica-

20:14 And Moses sent messengers from Kadesh unto the tus est in eis 20:14 misit interea nuntios Moses de Cades

king of Edom, Thus saith thy brother Israel, Thou knowest all the travail that hath befallen us:

ad regem Edom qui dicerent haec mandat frater tuus Israhel

20:15 How our fathers went down into Egypt, and we have nosti omnem laborem qui adprehendit nos 20:15 quomodo

dwelt in Egypt a long time; and the Egyptians vexed us, and our fathers:

descenderint patres nostri in Aegyptum et habitauerimus ibi multo tempore adflixerintque nos Aegyptii et patres nostros

20:16 And when we cried unto the LORD, he heard our 20:16 et quomodo clamauerimus ad Dominum et exaudierit

voice, and sent an angel, and hath brought us forth out of Egypt: and, behold, we are in Kadesh, a city in the

uttermost of thy border: nos miseritque angelum qui eduxerit nos de Aegypto ecce 20:17 Let us pass, I pray thee, through thy country: we in urbe Cades quae est in extremis finibus tuis positi 20:17

will not pass through the fields, or through the vineyards, neither will we drink of the water of the wells: we will go by the king’s high way, we will not turn to the right hand nor to the left, until we have passed thy borders.

obsecramus ut nobis transire liceat per terram tuam non ibimus per agros nec per uineas non bibemus aquas de puteis tuis sed gradiemur uia publica nec ad dextram nec ad si-

20:18 And Edom said unto him, Thou shalt not pass by nistram declinantes donec transeamus terminos tuos 20:18

me, lest I come out against thee with the sword.

cui respondit Edom non transibis per me alioquin armatus

20:19 And the children of Israel said unto him, We will go occurram tibi 20:19 dixeruntque filii Israhel per tritam gra-

by the high way: and if I and my cattle drink of thy water, then I will pay for it: I will only, without doing anything

else, go through on my feet. diemur uiam et si biberimus aquas tuas nos et pecora nostra

dabimus quod iustum est nulla erit in pretio difficultas tan-

20:20 And he said, Thou shalt not go through. And Edom tum uelociter transeamus 20:20 at ille respondit non trans-

came out against him with much people, and with a strong hand.

ibis statimque egressus est obuius cum infinita multitudine et

20:21 Thus Edom refused to give Israel passage through manu forti 20:21 nec uoluit adquiescere deprecanti ut con-

his border: wherefore Israel turned away from him.

cederet transitum per fines suos quam ob rem deuertit ab

20:22 And the children of Israel, even the whole congrega- eo Israhel 20:22 cumque castra mouissent de Cades uenetion, journeyed from Kadesh, and came unto mount Hor, by the coast of the land of Edom.

runt in montem Or qui est in finibus terrae Edom 20:23 ubi

20:23 And the LORD spake unto Moses and Aaron in mount Hor, saying, 20:24 Aaron shall be gathered unto his people: for he shall locutus est Dominus ad Mosen 20:24 pergat inquit Aaron not enter into the land which I have given unto the children of Israel, because ye rebelled against my word at the water

of Meribah. ad populos suos non enim intrabit terram quam dedi filiis

Israhel eo quod incredulus fuerit ori meo ad aquas Contra-

20:25 Take Aaron and Eleazar his son, and bring them up dictionis 20:25 tolle Aaron et filium eius cum eo et duces

unto mount Hor: 20:26 And strip Aaron of his garments, and put them upon eos in montem Or 20:26 cumque nudaueris patrem ueste Eleazar his son: and Aaron shall be gathered unto his people, and shall die there.

sua indues ea Eleazarum filium eius et Aaron colligetur et

20:27 And Moses did as the LORD commanded: and they morietur ibi 20:27 fecit Moses ut praeceperat Dominus et

went up into mount Hor in the sight of all the congregation.

ascenderunt in montem Or coram omni multitudine 20:28

20:28 And Moses stripped Aaron of his garments, and put them upon Eleazar his son;

cumque Aaron spoliasset uestibus suis induit eis Eleazarum

20:29 and Aaron died there in the top of the mount: and filium eius 20:29 illo mortuo in montis supercilio descen-

Moses and Eleazar came down from the mount.

dit cum Eleazaro 20:30 omnis autem multitudo uidens oc- 20:30 And when all the congregation saw that Aaron was

dead, they mourned for Aaron thirty days, even all the house of Israel.

cubuisse Aaron fleuit super eo triginta diebus per cunctas familias suas

21:1 quod cum audisset Chananeus rex Arad qui habitabat 21:1 And when king Arad the Canaanite, which dwelt in

the south, heard tell that Israel came by the way of the spies; then he fought against Israel, and took some of them

ad meridiem uenisse scilicet Israhel per exploratorum uiam prisoners. pugnauit contra illum et uictor existens duxit ex eo praedam 21:2 at Israhel uoto se Domino obligans ait si tradideris po- 21:2 And Israel vowed a vow unto the LORD, and said, If

thou wilt indeed deliver this people into my hand, then I will utterly destroy their cities.

pulum istum in manu mea delebo urbes eius 21:3 exaudi-

21:3 And the LORD hearkened to the voice of Israel, and delivered up the Canaanites; and they utterly destroyed them and their cities: and he called the name of the place Hormah.

uitque Dominus preces Israhel et tradidit Chananeum quem ille interfecit subuersis urbibus eius et uocauit nomen loci illius Horma id est anathema 21:4 profecti sunt autem et de 21:4 And they journeyed from mount Hor by the way of

the Red sea, to compass the land of Edom: and the soul of the people was much discouraged because of the way.

monte Or per uiam quae ducit ad mare Rubrum ut circumirent terram Edom et taedere coepit populum itineris ac laboris 21:5 locutusque contra Deum et Mosen ait cur eduxisti 21:5 And the people spake against God, and against

Moses, Wherefore have ye brought us up out of Egypt to die in the wilderness? for there is no bread, neither is there

nos de Aegypto ut moreremur in solitudine deest panis non any water; and our soul loatheth this light bread. sunt aquae anima nostra iam nausiat super cibo isto leuis-

simo 21:6 quam ob rem misit Dominus in populum igni- 21:6 And the LORD sent fiery serpents among the people,

and they bit the people; and much people of Israel died.

tos serpentes ad quorum plagas et mortes plurimorum 21:7 21:7 Therefore the people came to Moses, and said, We

have sinned, for we have spoken against the LORD, and against thee; pray unto the LORD, that he take away the

uenerunt ad Mosen atque dixerunt peccauimus quia locuti serpents from us. And Moses prayed for the people. sumus contra Dominum et te ora ut tollat a nobis serpentes orauit Moses pro populo 21:8 et locutus est Dominus ad 21:8 And the LORD said unto Moses, Make thee a fiery

serpent, and set it upon a pole: and it shall come to pass, that every one that is bitten, when he looketh upon it, shall

eum fac serpentem et pone eum pro signo qui percussus as- live. pexerit eum uiuet 21:9 fecit ergo Moses serpentem aeneum 21:9 And Moses made a serpent of brass, and put it upon

a pole, and it came to pass, that if a serpent had bitten any man, when he beheld the serpent of brass, he lived.

et posuit pro signo quem cum percussi aspicerent sanabantur 21:10 profectique filii Israhel castrametati sunt in Oboth 21:10 And the children of Israel set forward, and pitched

in Oboth.

21:11 unde egressi fixere tentoria in Hieabarim in solitudine 21:11 And they journeyed from Oboth, and pitched at Ije-

abarim, in the wilderness which is before Moab, toward the sunrising.

quae respicit Moab contra orientalem plagam 21:12 et inde

21:12 From thence they removed, and pitched in the valley of Zared.

mouentes uenerunt ad torrentem Zared 21:13 quem relin- 21:13 From thence they removed, and pitched on the other

side of Arnon, which is in the wilderness that cometh out of the coasts of the Amorites: for Arnon is the border of

quentes castrametati sunt contra Arnon quae est in deserto Moab, between Moab and the Amorites. et prominet in finibus Amorrei siquidem Arnon terminus est Moab diuidens Moabitas et Amorreos 21:14 unde dicitur 21:14 Wherefore it is said in the book of the wars of the

LORD, What he did in the Red sea, and in the brooks of Arnon,

in libro bellorum Domini sicut fecit in mari Rubro sic faciet in torrentibus Arnon 21:15 scopuli torrentium inclinati 21:15 And at the stream of the brooks that goeth down to

the dwelling of Ar, and lieth upon the border of Moab.

Biblia Sacra Vulgata

sunt ut requiescerent in Ar et recumberent in finibus Moa-

21:16 And from thence they went to Beer: that is the well bitarum 21:16 ex eo loco apparuit puteus super quo locutus

whereof the LORD spake unto Moses, Gather the people together, and I will give them water.

est Dominus ad Mosen congrega populum et dabo ei aquam

21:17 Then Israel sang this song, Spring up, O well; sing 21:17 tunc cecinit Israhel carmen istud ascendat puteus con-

ye unto it:

21:18 The princes digged the well, the nobles of the peo- cinebant 21:18 puteus quem foderunt principes et paraue-

ple digged it, by the direction of the lawgiver, with their staves. And from the wilderness they went to Mattanah:

runt duces multitudinis in datore legis et in baculis suis de

21:19 And from Mattanah to Nahaliel: and from Nahaliel solitudine Matthana 21:19 de Matthana Nahalihel de Naha-

to Bamoth:

21:20 And from Bamoth in the valley, that is in the coun- lihel in Bamoth 21:20 de Bamoth uallis est in regione Moab

try of Moab, to the top of Pisgah, which looketh toward Jeshimon.

in uertice Phasga et quod respicit contra desertum 21:21 mi-

21:21 And Israel sent messengers unto Sihon king of the Amorites, saying,

sit autem Israhel nuntios ad Seon regem Amorreorum dicens

21:22 Let me pass through thy land: we will not turn into 21:22 obsecro ut transire mihi liceat per terram tuam non de-

the fields, or into the vineyards; we will not drink of the waters of the well: but we will go along by the king’s high

way, until we be past thy borders. clinabimus in agros et uineas non bibemus aquas ex puteis 21:23 And Sihon would not suffer Israel to pass through uia regia gradiemur donec transeamus terminos tuos 21:23

his border: but Sihon gathered all his people together, and went out against Israel into the wilderness: and he came

to Jahaz, and fought against Israel. qui concedere noluit ut transiret Israhel per fines suos quin

potius exercitu congregato egressus est obuiam in desertum

21:24 And Israel smote him with the edge of the sword, et uenit in Iasa pugnauitque contra eum 21:24 a quo percus-

and possessed his land from Arnon unto Jabbok, even unto the children of Ammon: for the border of the children of

Ammon was strong. sus est in ore gladii et possessa est terra eius ab Arnon us-

que Iebboc et filios Ammon quia forti praesidio tenebantur

21:25 And Israel took all these cities: and Israel dwelt in termini Ammanitarum 21:25 tulit ergo Israhel omnes ciuiall the cities of the Amorites, in Heshbon, and in all the villages thereof.

tates eius et habitauit in urbibus Amorrei in Esebon scilicet

21:26 For Heshbon was the city of Sihon the king of et uiculis eius 21:26 urbs Esebon fuit regis Seon Amorrei

the Amorites, who had fought against the former king of Moab, and taken all his land out of his hand, even unto

Arnon. qui pugnauit contra regem Moab et tulit omnem terram quae 21:27 Wherefore they that speak in proverbs say, Come dicionis illius fuerat usque Arnon 21:27 idcirco dicitur in

into Heshbon, let the city of Sihon be built and prepared:

prouerbio uenite in Esebon aedificetur et construatur ciuitas

21:28 For there is a fire gone out of Heshbon, a flame from Seon 21:28 ignis egressus est de Esebon flamma de oppido

the city of Sihon: it hath consumed Ar of Moab, and the lords of the high places of Arnon.

Seon et deuorauit Ar Moabitarum et habitatores excelsorum

21:29 Woe to thee, Moab! thou art undone, O people of Arnon 21:29 uae tibi Moab peristi popule Chamos dedit fi-

Chemosh: he hath given his sons that escaped, and his daughters, into captivity unto Sihon king of the Amorites.

lios eius in fugam et filias in captiuitatem regi Amorreorum

21:30 We have shot at them; Heshbon is perished even Seon 21:30 iugum ipsorum disperiit ab Esebon usque Di-

unto Dibon, and we have laid them waste even unto Nophah, which reacheth unto Medeba.

bon lassi peruenerunt in Nophe et usque Medaba 21:31 ha-

21:31 Thus Israel dwelt in the land of the Amorites.

21:32 And Moses sent to spy out Jaazer, and they took bitauit itaque Israhel in terra Amorrei 21:32 misitque Mo-

the villages thereof, and drove out the Amorites that were there.

ses qui explorarent Iazer cuius ceperunt uiculos et posse-

21:33 And they turned and went up by the way of Bashan: derunt habitatores 21:33 uerteruntque se et ascenderunt per

and Og the king of Bashan went out against them, he, and all his people, to the battle at Edrei.

uiam Basan et occurrit eis Og rex Basan cum omni populo suo pugnaturus in Edrai 21:34 dixitque Dominus ad Mo- 21:34 And the LORD said unto Moses, Fear him not: for

I have delivered him into thy hand, and all his people, and his land; and thou shalt do to him as thou didst unto Sihon

sen ne timeas eum quia in manu tua tradidi illum et omnem king of the Amorites, which dwelt at Heshbon. populum ac terram eius faciesque illi sicut fecisti Seon regi Amorreorum habitatori Esebon 21:35 percusserunt igitur et 21:35 So they smote him, and his sons, and all his people,

until there was none left him alive: and they possessed his land.

hunc cum filiis suis uniuersumque populum eius usque ad internicionem et possederunt terram illius

22:1 profectique castrametati sunt in campestribus Moab 22:1 And the children of Israel set forward, and pitched in

the plains of Moab on this side Jordan by Jericho.

ubi trans Iordanem Hierichus sita est 22:2 uidens autem Ba- 22:2 And Balak the son of Zippor saw all that Israel had

done to the Amorites.

lac filius Sepphor omnia quae fecerat Israhel Amorreo 22:3 22:3 And Moab was sore afraid of the people, because

they were many: and Moab was distressed because of the children of Israel.

et quod pertimuissent eum Moabitae et impetum eius ferre non possent 22:4 dixit ad maiores natu Madian ita delebit 22:4 And Moab said unto the elders of Midian, Now shall

this company lick up all that are round about us, as the ox licketh up the grass of the field. And Balak the son of

hic populus omnes qui in nostris finibus commorantur quo- Zippor was king of the Moabites at that time. modo solet bos herbas usque ad radices carpere ipse erat eo

tempore rex in Moab 22:5 misit ergo nuntios ad Balaam 22:5 He sent messengers therefore unto Balaam the son of Beor to Pethor, which is by the river of the land of the children of his people, to call him, saying, Behold, there is a people come out from Egypt: behold, they cover the face of the earth, and they abide over against me:

filium Beor ariolum qui habitabat super flumen terrae filiorum Ammon ut uocarent eum et dicerent ecce egressus est populus ex Aegypto qui operuit superficiem terrae sedens contra me 22:6 ueni igitur et maledic populo huic quia for- 22:6 Come now therefore, I pray thee, curse me this peo-

ple; for they are too mighty for me: peradventure I shall prevail, that we may smite them, and that I may drive them out of the land: for I wot that he whom thou blessest is blessed, and he whom thou cursest is cursed.

tior me est si quo modo possim percutere et eicere eum de terra mea noui enim quod benedictus sit cui benedixeris et maledictus in quem maledicta congesseris 22:7 perrexerunt 22:7 And the elders of Moab and the elders of Midian de-

parted with the rewards of divination in their hand; and they came unto Balaam, and spake unto him the words of

seniores Moab et maiores natu Madian habentes diuinationis Balak. pretium in manibus cumque uenissent ad Balaam et narrassent ei omnia uerba Balac 22:8 ille respondit manete hic 22:8 And he said unto them, Lodge here this night, and I

will bring you word again, as the LORD shall speak unto me: and the princes of Moab abode with Balaam.

nocte et respondebo quicquid mihi dixerit Dominus manentibus illis apud Balaam uenit Deus et ait ad eum 22:9 quid 22:9 And God came unto Balaam, and said, What men are

these with thee?

sibi uolunt homines isti apud te 22:10 respondit Balac filius 22:10 And Balaam said unto God, Balak the son of Zippor,

king of Moab, hath sent unto me, saying,

Sepphor rex Moabitarum misit ad me 22:11 dicens ecce po- 22:11 Behold, there is a people come out of Egypt, which

covereth the face of the earth: come now, curse me them; peradventure I shall be able to overcome them, and drive

pulus qui egressus est de Aegypto operuit superficiem terrae them out. ueni et maledic ei si quo modo possim pugnans abicere eum 22:12 dixitque Deus ad Balaam noli ire cum eis neque ma- 22:12 And God said unto Balaam, Thou shalt not go with

them; thou shalt not curse the people: for they are blessed.

ledicas populo quia benedictus est 22:13 qui mane consur- 22:13 And Balaam rose up in the morning, and said unto

the princes of Balak, Get you into your land: for the LORD refuseth to give me leave to go with you.

Biblia Sacra Vulgata

gens dixit ad principes ite in terram uestram quia prohibuit

22:14 And the princes of Moab rose up, and they went me Deus uenire uobiscum 22:14 reuersi principes dixerunt

unto Balak, and said, Balaam refuseth to come with us.

22:15 And Balak sent yet again princes, more, and more ad Balac noluit Balaam uenire nobiscum 22:15 rursum ille

honourable than they.

22:16 And they came to Balaam, and said to him, Thus multo plures et nobiliores quam ante miserat misit 22:16

saith Balak the son of Zippor, Let nothing, I pray thee, hinder thee from coming unto me:

qui cum uenissent ad Balaam dixerunt sic dicit Balac fi-

22:17 For I will promote thee unto very great honour, and lius Sepphor ne cuncteris uenire ad me 22:17 paratum ho-I will do whatsoever thou sayest unto me: come therefore, I pray thee, curse me this people.

norare te et quicquid uolueris dare ueni et maledic populo

22:18 And Balaam answered and said unto the servants of isti 22:18 respondit Balaam si dederit mihi Balac plenam

Balak, If Balak would give me his house full of silver and gold, I cannot go beyond the word of the LORD my God,

to do less or more. domum suam argenti et auri non potero inmutare uerbum 22:19 Now therefore, I pray you, tarry ye also here this Domini Dei mei ut uel plus uel minus loquar 22:19 obsecro

night, that I may know what the LORD will say unto me more.

ut hic maneatis etiam hac nocte et scire queam quid mihi rur-

22:20 And God came unto Balaam at night, and said unto sum respondeat Dominus 22:20 uenit ergo Deus ad Balaam

him, If the men come to call thee, rise up, and go with them; but yet the word which I shall say unto thee, that

shalt thou do. nocte et ait ei si uocare te uenerunt homines isti surge et uade 22:21 And Balaam rose up in the morning, and saddled his cum eis ita dumtaxat ut quod tibi praecepero facias 22:21

ass, and went with the princes of Moab.

surrexit Balaam mane et strata asina profectus est cum eis

22:22 And God’s anger was kindled because he went: and 22:22 et iratus est Deus stetitque angelus Domini in uia con-

the angel of the LORD stood in the way for an adversary against him. Now he was riding upon his ass, and his two

servants were with him. tra Balaam qui sedebat asinae et duos pueros habebat secum 22:23 And the ass saw the angel of the LORD standing 22:23 cernens asina angelum stantem in uia euaginato gladio

in the way, and his sword drawn in his hand: and the ass turned aside out of the way, and went into the field: and

Balaam smote the ass, to turn her into the way. auertit se de itinere et ibat per agrum quam cum uerberaret 22:24 But the angel of the LORD stood in a path of the Balaam et uellet ad semitam reducere 22:24 stetit angelus

vineyards, a wall being on this side, and a wall on that side.

in angustiis duarum maceriarum quibus uineae cingebantur

22:25 And when the ass saw the angel of the LORD, she 22:25 quem uidens asina iunxit se parieti et adtriuit sedentis

thrust herself unto the wall, and crushed Balaam’s foot against the wall: and he smote her again.

pedem at ille iterum uerberabat 22:26 et nihilominus ange-

22:26 And the angel of the LORD went further, and stood in a narrow place, where was no way to turn either to the right hand or to the left.

lus ad locum angustum transiens ubi nec ad dextram nec ad

22:27 And when the ass saw the angel of the LORD, she sinistram poterat deuiari obuius stetit 22:27 cumque uidis-

fell down under Balaam: and Balaam’s anger was kindled, and he smote the ass with a staff.

set asina stantem angelum concidit sub pedibus sedentis qui

22:28 And the LORD opened the mouth of the ass, and iratus uehementius caedebat fuste latera 22:28 aperuitque

she said unto Balaam, What have I done unto thee, that thou hast smitten me these three times?

Dominus os asinae et locuta est quid feci tibi cur percutis me

22:29 And Balaam said unto the ass, Because thou hast ecce iam tertio 22:29 respondit Balaam quia commeruisti

mocked me: I would there were a sword in mine hand, for now would I kill thee.

et inlusisti mihi utinam haberem gladium ut te percuterem

22:30 And the ass said unto Balaam, Am not I thine ass, 22:30 dixit asina nonne animal tuum sum cui semper sedere

upon which thou hast ridden ever since I was thine unto this day? was I ever wont to do so unto thee? And he said,

Nay. consuesti usque in praesentem diem dic quid simile umquam 22:31 Then the LORD opened the eyes of Balaam, and he fecerim tibi at ille ait numquam 22:31 protinus aperuit Do-

saw the angel of the LORD standing in the way, and his sword drawn in his hand: and he bowed down his head, and fell flat on his face.

minus oculos Balaam et uidit angelum stantem in uia euaginato gladio adorauitque eum pronus in terram 22:32 cui 22:32 And the angel of the LORD said unto him, Where-

fore hast thou smitten thine ass these three times? behold, I went out to withstand thee, because thy way is perverse

angelus cur inquit tertio uerberas asinam tuam ego ueni ut before me: aduersarer tibi quia peruersa est uia tua mihique contraria 22:33 et nisi asina declinasset de uia dans locum resistenti 22:33 And the ass saw me, and turned from me these three

times: unless she had turned from me, surely now also I had slain thee, and saved her alive.

te occidissem et illa uiueret 22:34 dixit Balaam peccaui ne-

22:34 And Balaam said unto the angel of the LORD, I have sinned; for I knew not that thou stoodest in the way against me: now therefore, if it displease thee, I will get me back again.

sciens quod tu stares contra me et nunc si displicet tibi ut uadam reuertar 22:35 ait angelus uade cum istis et caue ne 22:35 And the angel of the LORD said unto Balaam, Go

with the men: but only the word that I shall speak unto thee, that thou shalt speak. So Balaam went with the

aliud quam praecepero tibi loquaris iuit igitur cum principi- princes of Balak. bus 22:36 quod cum audisset Balac egressus est in occur- 22:36 And when Balak heard that Balaam was come, he

went out to meet him unto a city of Moab, which is in the border of Arnon, which is in the utmost coast.

sum eius in oppido Moabitarum quod situm est in extremis finibus Arnon 22:37 dixitque ad Balaam misi nuntios ut uo- 22:37 And Balak said unto Balaam, Did I not earnestly

send unto thee to call thee? wherefore camest thou not unto me? am I not able indeed to promote thee to honour?

carent te cur non statim uenisti ad me an quia mercedem aduentui tuo reddere nequeo 22:38 cui ille respondit ecce ad- 22:38 And Balaam said unto Balak, Lo, I am come unto

thee: have I now any power at all to say any thing? the word that God putteth in my mouth, that shall I speak.

sum numquid loqui potero aliud nisi quod Deus posuerit in ore meo 22:39 perrexerunt ergo simul et uenerunt in urbem 22:39 And Balaam went with Balak, and they came unto

Kirjathhuzoth.

quae in extremis regni eius finibus erat 22:40 cumque occi- 22:40 And Balak offered oxen and sheep, and sent to Bal-

aam, and to the princes that were with him.

disset Balac boues et oues misit ad Balaam et principes qui cum eo erant munera 22:41 mane autem facto duxit eum ad 22:41 And it came to pass on the morrow, that Balak took

Balaam, and brought him up into the high places of Baal, that thence he might see the utmost part of the people.

excelsa Baal et intuitus est extremam partem populi

23:1 dixitque Balaam ad Balac aedifica mihi hic septem 23:1 And Balaam said unto Balak, Build me here seven

altars, and prepare me here seven oxen and seven rams.

aras et para totidem uitulos eiusdemque numeri arietes 23:2 23:2 And Balak did as Balaam had spoken; and Balak and

Balaam offered on every altar a bullock and a ram.

cumque fecisset iuxta sermonem Balaam inposuerunt simul uitulum et arietem super aram 23:3 dixitque Balaam ad Ba- 23:3 And Balaam said unto Balak, Stand by thy burnt of-

fering, and I will go: peradventure the LORD will come to meet me: and whatsoever he sheweth me I will tell thee.

lac sta paulisper iuxta holocaustum tuum donec uadam si And he went to an high place. forte occurrat mihi Dominus et quodcumque imperauerit loquar tibi 23:4 cumque abisset uelociter occurrit ei Deus lo- 23:4 And God met Balaam: and he said unto him, I have

prepared seven altars, and I have offered upon every altar a bullock and a ram.

cutusque ad eum Balaam septem inquit aras erexi et inposui uitulum et arietem desuper 23:5 Dominus autem posuit 23:5 And the LORD put a word in Balaam’s mouth, and

said, Return unto Balak, and thus thou shalt speak.

uerbum in ore eius et ait reuertere ad Balac et haec loqueris 23:6 reuersus inuenit stantem Balac iuxta holocaustum 23:6 And he returned unto him, and, lo, he stood by his

burnt sacrifice, he, and all the princes of Moab.

suum et omnes principes Moabitarum 23:7 adsumptaque 23:7 And he took up his parable, and said, Balak the king of Moab hath brought me from Aram, out of the mountains of the east, saying, Come, curse me Jacob, and come, defy

parabola sua dixit de Aram adduxit me Balac rex Moabi- Israel.

Biblia Sacra Vulgata

tarum de montibus orientis ueni inquit et maledic Iacob pro-

23:8 How shall I curse, whom God hath not cursed? or pera et detestare Israhel 23:8 quomodo maledicam cui non

how shall I defy, whom the LORD hath not defied?

maledixit Deus qua ratione detester quem Dominus non de-

23:9 For from the top of the rocks I see him, and from the testatur 23:9 de summis silicibus uidebo eum et de collibus

hills I behold him: lo, the people shall dwell alone, and shall not be reckoned among the nations.

considerabo illum populus solus habitabit et inter gentes non

23:10 Who can count the dust of Jacob, and the number reputabitur 23:10 quis dinumerare possit puluerem Iacob et

of the fourth part of Israel? Let me die the death of the righteous, and let my last end be like his!

nosse numerum stirpis Israhel moriatur anima mea morte iu-

23:11 And Balak said unto Balaam, What hast thou done storum et fiant nouissima mea horum similia 23:11 dixitque

unto me? I took thee to curse mine enemies, and, behold, thou hast blessed them altogether.

Balac ad Balaam quid est hoc quod agis ut malediceres in-

23:12 And he answered and said, Must I not take heed to imicis uocaui te et tu e contrario benedicis eis 23:12 cui ille

speak that which the LORD hath put in my mouth?

respondit num aliud possum loqui nisi quod iusserit Domi-

23:13 And Balak said unto him, Come, I pray thee, with nus 23:13 dixit ergo Balac ueni mecum in alterum locum

me unto another place, from whence thou mayest see them: thou shalt see but the utmost part of them, and shalt

not see them all: and curse me them from thence. unde partem Israhelis uideas et totum uidere non possis inde 23:14 And he brought him into the field of Zophim, to the maledicito ei 23:14 cumque duxisset eum in locum subli-

top of Pisgah, and built seven altars, and offered a bullock and a ram on every altar.

mem super uerticem montis Phasga aedificauit Balaam sep-

23:15 And he said unto Balak, Stand here by thy burnt tem aras et inpositis supra uitulo atque ariete 23:15 dixit ad

offering, while I meet the LORD yonder.

Balac sta hic iuxta holocaustum tuum donec ego pergam ob-

23:16 And the LORD met Balaam, and put a word in his uius 23:16 cui cum Dominus occurrisset posuissetque uer-

mouth, and said, Go again unto Balak, and say thus.

bum in ore eius ait reuertere ad Balac et haec loqueris ei

23:17 And when he came to him, behold, he stood by his 23:17 reuersus inuenit eum stantem iuxta holocaustum suum

burnt offering, and the princes of Moab with him. And Balak said unto him, What hath the LORD spoken?

et principes Moabitarum cum eo ad quem Balac quid inquit

23:18 And he took up his parable, and said, Rise up, Balak, locutus est Dominus 23:18 at ille adsumpta parabola sua ait

and hear; hearken unto me, thou son of Zippor: 23:19 God is not a man, that he should lie; neither the son sta Balac et ausculta audi fili Sepphor 23:19 non est Deof man, that he should repent: hath he said, and shall he not do it? or hath he spoken, and shall he not make it

good? us quasi homo ut mentiatur nec ut filius hominis ut mutetur 23:20 Behold, I have received commandment to bless: and dixit ergo et non faciet locutus est et non implebit 23:20 ad

he hath blessed; and I cannot reverse it.

benedicendum adductus sum benedictionem prohibere non

23:21 He hath not beheld iniquity in Jacob, neither hath ualeo 23:21 non est idolum in Iacob nec uidetur simulache seen perverseness in Israel: the LORD his God is with him, and the shout of a king is among them.

rum in Israhel Dominus Deus eius cum eo est et clangor

23:22 God brought them out of Egypt; he hath as it were uictoriae regis in illo 23:22 Deus eduxit eum de Aegypto

the strength of an unicorn.

23:23 Surely there is no enchantment against Jacob, nei- cuius fortitudo similis est rinocerotis 23:23 non est augu-

ther is there any divination against Israel: according to this time it shall be said of Jacob and of Israel, What hath God

wrought! rium in Iacob nec diuinatio in Israhel temporibus suis dicetur 23:24 Behold, the people shall rise up as a great lion, and Iacob et Israheli quid operatus sit Deus 23:24 ecce populus

lift up himself as a young lion: he shall not lie down until he eat of the prey, and drink the blood of the slain.

ut leaena consurget et quasi leo erigetur non accubabit donec deuoret praedam et occisorum sanguinem bibat 23:25 dixit- 23:25 And Balak said unto Balaam, Neither curse them at

all, nor bless them at all.

que Balac ad Balaam nec maledicas ei nec benedicas 23:26 23:26 But Balaam answered and said unto Balak, Told not

I thee, saying, All that the LORD speaketh, that I must do?

et ille nonne ait dixi tibi quod quicquid mihi Deus imperaret hoc facerem 23:27 et ait Balac ad eum ueni et ducam te 23:27 And Balak said unto Balaam, Come, I pray thee,

I will bring thee unto another place; peradventure it will please God that thou mayest curse me them from thence.

ad alium locum si forte placeat Deo ut inde maledicas eis 23:28 cumque duxisset eum super uerticem montis Phogor 23:28 And Balak brought Balaam unto the top of Peor, that

looketh toward Jeshimon.

qui respicit solitudinem 23:29 dixit ei Balaam aedifica mihi 23:29 And Balaam said unto Balak, Build me here seven

altars, and prepare me here seven bullocks and seven rams.

hic septem aras et para totidem uitulos eiusdemque numeri arietes 23:30 fecit Balac ut Balaam dixerat inposuitque ui- 23:30 And Balak did as Balaam had said, and offered a

bullock and a ram on every altar.

tulos et arietes per singulas aras

24:1 cumque uidisset Balaam quod placeret Domino ut be- 24:1 And when Balaam saw that it pleased the LORD to

bless Israel, he went not, as at other times, to seek for enchantments, but he set his face toward the wilderness.

nediceret Israheli nequaquam abiit ut ante perrexerat ut augurium quaereret sed dirigens contra desertum uultum suum 24:2 et eleuans oculos uidit Israhel in tentoriis commorantem 24:2 And Balaam lifted up his eyes, and he saw Israel

abiding in his tents according to their tribes; and the spirit of God came upon him.

per tribus suas et inruente in se spiritu Dei 24:3 adsumpta

24:3 And he took up his parable, and said, Balaam the son of Beor hath said, and the man whose eyes are open hath said:

parabola ait dixit Balaam filius Beor dixit homo cuius obturatus est oculus 24:4 dixit auditor sermonum Dei qui ui- 24:4 He hath said, which heard the words of God, which

saw the vision of the Almighty, falling into a trance, but having his eyes open:

sionem Omnipotentis intuitus est qui cadit et sic aperiuntur oculi eius 24:5 quam pulchra tabernacula tua Iacob et ten- 24:5 How goodly are thy tents, O Jacob, and thy taberna-

cles, O Israel!

toria tua Israhel 24:6 ut ualles nemorosae ut horti iuxta flu- 24:6 As the valleys are they spread forth, as gardens by

the river’s side, as the trees of lign aloes which the LORD hath planted, and as cedar trees beside the waters.

uios inrigui ut tabernacula quae fixit Dominus quasi cedri propter aquas 24:7 fluet aqua de situla eius et semen il- 24:7 He shall pour the water out of his buckets, and his

seed shall be in many waters, and his king shall be higher than Agag, and his kingdom shall be exalted.

lius erit in aquas multas tolletur propter Agag rex eius et auferetur regnum illius 24:8 Deus eduxit illum de Aegypto 24:8 God brought him forth out of Egypt; he hath as it

were the strength of an unicorn: he shall eat up the nations his enemies, and shall break their bones, and pierce them

cuius fortitudo similis est rinocerotis deuorabunt gentes ho- through with his arrows. stes illius ossaque eorum confringent et perforabunt sagittis 24:9 accubans dormiuit ut leo et quasi leaena quam susci- 24:9 He couched, he lay down as a lion, and as a great

lion: who shall stir him up? Blessed is he that blesseth thee, and cursed is he that curseth thee.

tare nullus audebit qui benedixerit tibi erit ipse benedictus qui maledixerit in maledictione reputabitur 24:10 iratusque 24:10 And Balak’s anger was kindled against Balaam, and

he smote his hands together: and Balak said unto Balaam, I called thee to curse mine enemies, and, behold, thou hast

Balac contra Balaam conplosis manibus ait ad maledicen- altogether blessed them these three times. dum inimicis meis uocaui te quibus e contrario tertio benedixisti 24:11 reuertere ad locum tuum decreueram qui- 24:11 Therefore now flee thou to thy place: I thought to

promote thee unto great honour; but, lo, the LORD hath kept thee back from honour.

dem magnifice honorare te sed Dominus priuauit te honore

Biblia Sacra Vulgata

24:12 And Balaam said unto Balak, Spake I not also to thy disposito 24:12 respondit Balaam ad Balac nonne nuntiis

messengers which thou sentest unto me, saying, 24:13 If Balak would give me his house full of silver and tuis quos misisti ad me dixi 24:13 si dederit mihi Balac gold, I cannot go beyond the commandment of the LORD, to do either good or bad of mine own mind; but what the

LORD saith, that will I speak? plenam domum suam argenti et auri non potero praeterire sermonem Domini Dei mei ut uel boni quid uel mali proferam ex corde meo sed quicquid Dominus dixerit hoc lo-

24:14 And now, behold, I go unto my people: come there- quar 24:14 uerumtamen pergens ad populum meum dabo fore, and I will advertise thee what this people shall do to thy people in the latter days.

consilium quid populus tuus huic populo faciat extremo tem-

24:15 And he took up his parable, and said, Balaam the pore 24:15 sumpta igitur parabola rursum ait dixit Balaam

son of Beor hath said, and the man whose eyes are open hath said:

filius Beor dixit homo cuius obturatus est oculus 24:16 dixit

24:16 He hath said, which heard the words of God, and knew the knowledge of the most High, which saw the vision of the Almighty, falling into a trance, but having his eyes open:

auditor sermonum Dei qui nouit doctrinam Altissimi et uisiones Omnipotentis uidet qui cadens apertos habet oculos

24:17 I shall see him, but not now: I shall behold him, 24:17 uidebo eum sed non modo intuebor illum sed non pro-

but not nigh: there shall come a Star out of Jacob, and a Sceptre shall rise out of Israel, and shall smite the corners

of Moab, and destroy all the children of Sheth. pe orietur stella ex Iacob et consurget uirga de Israhel et per-24:18 And Edom shall be a possession, Seir also shall be cutiet duces Moab uastabitque omnes filios Seth 24:18 et

a possession for his enemies; and Israel shall do valiantly.

erit Idumea possessio eius hereditas Seir cedet inimicis suis

24:19 Out of Jacob shall come he that shall have dominion, Israhel uero fortiter aget 24:19 de Iacob erit qui domine-

and shall destroy him that remaineth of the city.

24:20 And when he looked on Amalek, he took up his tur et perdat reliquias ciuitatis 24:20 cumque uidisset Ama-

parable, and said, Amalek was the first of the nations; but his latter end shall be that he perish for ever.

lech adsumens parabolam ait principium gentium Amalech

24:21 And he looked on the Kenites, and took up his para- cuius extrema perdentur 24:21 uidit quoque Cineum et ad-

ble, and said, Strong is thy dwellingplace, and thou puttest thy nest in a rock.

sumpta parabola ait robustum est quidem habitaculum tuum

24:22 Nevertheless the Kenite shall be wasted, until sed si in petra posueris nidum tuum 24:22 et fueris electus

Asshur shall carry thee away captive.

de stirpe Cain quamdiu poteris permanere Assur enim capiet

24:23 And he took up his parable, and said, Alas, who te 24:23 adsumptaque parabola iterum locutus est heu quis

shall live when God doeth this!

24:24 And ships shall come from the coast of Chittim, and uicturus est quando ista faciet Deus 24:24 uenient in trieri-

shall afflict Asshur, and shall afflict Eber, and he also shall perish for ever.

bus de Italia superabunt Assyrios uastabuntque Hebraeos et

24:25 And Balaam rose up, and went and returned to his ad extremum etiam ipsi peribunt 24:25 surrexitque Balaam

place: and Balak also went his way.

et reuersus est in locum suum Balac quoque uia qua uenerat rediit

25:1 And Israel abode in Shittim, and the people began to 25:1 morabatur autem eo tempore Israhel in Setthim et for-

commit whoredom with the daughters of Moab. 25:2 And they called the people unto the sacrifices of their nicatus est populus cum filiabus Moab 25:2 quae uocauegods: and the people did eat, and bowed down to their gods.

runt eos ad sacrificia sua at illi comederunt et adorauerunt

25:3 And Israel joined himself unto Baalpeor: and the deos earum 25:3 initiatusque est Israhel Beelphegor et ira-

anger of the LORD was kindled against Israel.

25:4 And the LORD said unto Moses, Take all the heads tus Dominus 25:4 ait ad Mosen tolle cunctos principes po-

of the people, and hang them up before the LORD against the sun, that the fierce anger of the LORD may be turned away from Israel.

puli et suspende eos contra solem in patibulis ut auertatur furor meus ab Israhel 25:5 dixitque Moses ad iudices Isra- 25:5 And Moses said unto the judges of Israel, Slay ye

every one his men that were joined unto Baalpeor.

hel occidat unusquisque proximos suos qui initiati sunt Beelphegor 25:6 et ecce unus de filiis Israhel intrauit coram 25:6 And, behold, one of the children of Israel came and

brought unto his brethren a Midianitish woman in the sight of Moses, and in the sight of all the congregation of the children of Israel, who were weeping before the door of the tabernacle of the congregation.

fratribus suis ad scortum madianitin uidente Mose et omni turba filiorum Israhel qui flebant ante fores tabernaculi 25:7 25:7 And when Phinehas, the son of Eleazar, the son of

Aaron the priest, saw it, he rose up from among the congregation, and took a javelin in his hand;

quod cum uidisset Finees filius Eleazari filii Aaron sacerdotis surrexit de medio multitudinis et arrepto pugione 25:8 25:8 And he went after the man of Israel into the tent, and

thrust both of them through, the man of Israel, and the woman through her belly. So the plague was stayed from

ingressus est post uirum israhelitem in lupanar et perfodit the children of Israel. ambos simul uirum scilicet et mulierem in locis genitalibus cessauitque plaga a filiis Israhel 25:9 et occisi sunt uiginti 25:9 And those that died in the plague were twenty and

four thousand.

quattuor milia homines 25:10 dixitque Dominus ad Mosen 25:10 And the LORD spake unto Moses, saying, 25:11 Finees filius Eleazari filii Aaron sacerdotis auertit iram 25:11 Phinehas, the son of Eleazar, the son of Aaron the

priest, hath turned my wrath away from the children of Israel, while he was zealous for my sake among them, that

meam a filiis Israhel quia zelo meo commotus est contra eos I consumed not the children of Israel in my jealousy. ut non ipse delerem filios Israhel in zelo meo 25:12 idcirco 25:12 Wherefore say, Behold, I give unto him my

covenant of peace:

loquere ad eos ecce do ei pacem foederis mei 25:13 et erit 25:13 And he shall have it, and his seed after him, even

the covenant of an everlasting priesthood; because he was zealous for his God, and made an atonement for the chil-

tam ipsi quam semini illius pactum sacerdotii sempiternum dren of Israel. quia zelatus est pro Deo suo et expiauit scelus filiorum Israhel 25:14 erat autem nomen uiri israhelitae qui occisus 25:14 Now the name of the Israelite that was slain, even

that was slain with the Midianitish woman, was Zimri, the son of Salu, a prince of a chief house among the Sime-

est cum Madianitide Zambri filius Salu dux de cognatione et onites. tribu Symeonis 25:15 porro mulier madianitis quae pariter 25:15 And the name of the Midianitish woman that was

slain was Cozbi, the daughter of Zur; he was head over a people, and of a chief house in Midian.

interfecta est uocabatur Chozbi filia Sur principis nobilissimi Madianitarum 25:16 locutusque est Dominus ad Mosen di- 25:16 And the LORD spake unto Moses, saying, cens 25:17 hostes uos sentiant Madianitae et percutite eos 25:17 Vex the Midianites, and smite them: 25:18 quia et ipsi hostiliter egerunt contra uos et decepere 25:18 For they vex you with their wiles, wherewith they

have beguiled you in the matter of Peor, and in the matter of Cozbi, the daughter of a prince of Midian, their sister,

insidiis per idolum Phogor et Chozbi filiam ducis Madian which was slain in the day of the plague for Peor’s sake. sororem suam quae percussa est in die plagae pro sacrilegio Phogor

26:1 postquam noxiorum sanguis effusus est dixit Domi- 26:1 And it came to pass after the plague, that the LORD

spake unto Moses and unto Eleazar the son of Aaron the priest, saying,

nus ad Mosen et Eleazarum filium Aaron sacerdotem 26:2

26:2 Take the sum of all the congregation of the children of Israel, from twenty years old and upward, throughout their fathers’ house, all that are able to go to war in Israel.

numerate omnem summam filiorum Israhel a uiginti annis et supra per domos et cognationes suas cunctos qui possunt ad bella procedere 26:3 locuti sunt itaque Moses et Elea- 26:3 And Moses and Eleazar the priest spake with them in

the plains of Moab by Jordan near Jericho, saying,

zar sacerdos in campestribus Moab super Iordanem contra

26:4 Take the sum of the people, from twenty years old Hierichum ad eos qui erant 26:4 a uiginti annis et supra

and upward; as the LORD commanded Moses and the children of Israel, which went forth out of the land of

Egypt. sicut Dominus imperarat quorum iste est numerus 26:5 Ru-

26:5 Reuben, the eldest son of Israel: the children of Reuben; Hanoch, of whom cometh the family of the Hanochites: of Pallu, the family of the Palluites:

ben primogenitus Israhel huius filius Enoch a quo familia

26:6 Of Hezron, the family of the Hezronites: of Carmi, Enochitarum et Phallu a quo familia Phalluitarum 26:6 et

the family of the Carmites.

Esrom a quo familia Esromitarum et Charmi a quo fami-

26:7 These are the families of the Reubenites: and they lia Charmitarum 26:7 hae sunt familiae de stirpe Ruben that were numbered of them were forty and three thousand and seven hundred and thirty.

quarum numerus inuentus est quadraginta tria milia et se-

26:8 And the sons of Pallu; Eliab. ptingenti triginta 26:8 filius Phallu Heliab 26:9 huius fi-

26:9 And the sons of Eliab; Nemuel, and Dathan, and Abiram. This is that Dathan and Abiram, which were famous in the congregation, who strove against Moses and against Aaron in the company of Korah, when they strove against the LORD:

lii Namuhel et Dathan et Abiram isti sunt Dathan et Abiram principes populi qui surrexerunt contra Mosen et Aaron in seditione Core quando aduersum Dominum rebella-

26:10 And the earth opened her mouth, and swallowed uerunt 26:10 et aperiens terra os suum deuorauit Core mo-

them up together with Korah, when that company died, what time the fire devoured two hundred and fifty men:

and they became a sign. rientibus plurimis quando conbusit ignis ducentos quinqua-26:11 Notwithstanding the children of Korah died not. ginta uiros et factum est grande miraculum 26:11 ut Core 26:12 The sons of Simeon after their families: of Nemuel, pereunte filii illius non perirent 26:12 filii Symeon per cog-

the family of the Nemuelites: of Jamin, the family of the Jaminites: of Jachin, the family of the Jachinites:

nationes suas Namuhel ab hoc familia Namuhelitarum Iamin ab hoc familia Iaminitarum Iachin ab hoc familia Iachinita-

26:13 Of Zerah, the family of the Zarhites: of Shaul, the rum 26:13 Zare ab hoc familia Zareitarum Saul ab hoc fafamily of the Shaulites.

26:14 These are the families of the Simeonites, twenty and milia Saulitarum 26:14 hae sunt familiae de stirpe Symeon two thousand and two hundred.

quarum omnis numerus fuit uiginti duo milia ducentorum

26:15 The children of Gad after their families: of Zephon, 26:15 filii Gad per cognationes suas Sephon ab hoc famithe family of the Zephonites: of Haggi, the family of the Haggites: of Shuni, the family of the Shunites:

lia Sephonitarum Aggi ab hoc familia Aggitarum Suni ab

26:16 Of Ozni, the family of the Oznites: of Eri, the family hoc familia Sunitarum 26:16 Ozni ab hoc familia Oznitaof the Erites:

26:17 Of Arod, the family of the Arodites: of Areli, the rum Heri ab hoc familia Heritarum 26:17 Arod ab hoc fafamily of the Arelites.

26:18 These are the families of the children of Gad accord- milia Aroditarum Arihel ab hoc familia Arihelitarum 26:18

ing to those that were numbered of them, forty thousand and five hundred.

istae sunt familiae Gad quarum omnis numerus fuit quadra-

26:19 The sons of Judah were Er and Onan: and Er and ginta milia quingentorum 26:19 filii Iuda Her et Onan qui

Onan died in the land of Canaan.

26:20 And the sons of Judah after their families were; of ambo mortui sunt in terra Chanaan 26:20 fueruntque filii Shelah, the family of the Shelanites: of Pharez, the family of the Pharzites: of Zerah, the family of the Zarhites.

Iuda per cognationes suas Sela a quo familia Selanitarum Phares a quo familia Pharesitarum Zare a quo familia Za-

26:21 And the sons of Pharez were; of Hezron, the family reitarum 26:21 porro filii Phares Esrom a quo familia Es-

of the Hezronites: of Hamul, the family of the Hamulites.

26:22 These are the families of Judah according to those romitarum et Amul a quo familia Amulitarum 26:22 istae that were numbered of them, threescore and sixteen thousand and five hundred.

sunt familiae Iuda quarum omnis numerus fuit septuaginta milia quingentorum 26:23 filii Isachar per cognationes suas 26:23 Of the sons of Issachar after their families: of

Tola, the family of the Tolaites: of Pua, the family of the Punites:

Thola a quo familia Tholaitarum Phua a quo familia Phuai-

tarum 26:24 Iasub a quo familia Iasubitarum Semran a quo 26:24 Of Jashub, the family of the Jashubites: of Shimron,

the family of the Shimronites.

familia Semranitarum 26:25 hae sunt cognationes Isachar 26:25 These are the families of Issachar according to those that were numbered of them, threescore and four thousand and three hundred.

quarum numerus fuit sexaginta quattuor milia trecentorum 26:26 filii Zabulon per cognationes suas Sared a quo familia 26:26 Of the sons of Zebulun after their families: of Sered,

the family of the Sardites: of Elon, the family of the Elonites: of Jahleel, the family of the Jahleelites.

Sareditarum Helon a quo familia Helonitarum Ialel a quo familia Ialelitarum 26:27 hae sunt cognationes Zabulon qua- 26:27 These are the families of the Zebulunites according

to those that were numbered of them, threescore thousand and five hundred.

rum numerus fuit sexaginta milia quingentorum 26:28 filii

26:28 The sons of Joseph after their families were Manasseh and Ephraim.

Ioseph per cognationes suas Manasse et Ephraim 26:29 de 26:29 Of the sons of Manasseh: of Machir, the family of

the Machirites: and Machir begat Gilead: of Gilead come the family of the Gileadites.

Manasse ortus est Machir a quo familia Machiritarum Machir genuit Galaad a quo familia Galaaditarum 26:30 Ga- 26:30 These are the sons of Gilead: of Jeezer, the family

of the Jeezerites: of Helek, the family of the Helekites:

laad habuit filios Hiezer a quo familia Hiezeritarum et Elec a quo familia Elecarum 26:31 et Asrihel a quo familia Asri- 26:31 And of Asriel, the family of the Asrielites: and of

Shechem, the family of the Shechemites:

helitarum et Sechem a quo familia Sechemitarum 26:32 et 26:32 And of Shemida, the family of the Shemidaites: and

of Hepher, the family of the Hepherites.

Semida a quo familia Semidatarum et Epher a quo familia Epheritarum 26:33 fuit autem Epher pater Salphaad qui fi- 26:33 And Zelophehad the son of Hepher had no sons, but

daughters: and the names of the daughters of Zelophehad were Mahlah, and Noah, Hoglah, Milcah, and Tirzah.

lios non habebat sed tantum filias quarum ista sunt nomina

Maala et Noa et Egla et Melcha et Thersa 26:34 hae sunt 26:34 These are the families of Manasseh, and those that were numbered of them, fifty and two thousand and seven hundred.

familiae Manasse et numerus earum quinquaginta duo milia septingentorum 26:35 filii autem Ephraim per cognationes 26:35 These are the sons of Ephraim after their families:

of Shuthelah, the family of the Shuthalhites: of Becher, the family of the Bachrites: of Tahan, the family of the

suas fuerunt hii Suthala a quo familia Suthalitarum Becher a Tahanites. quo familia Becheritarum Tehen a quo familia Tehenitarum 26:36 porro filius Suthala fuit Heran a quo familia Heranita- 26:36 And these are the sons of Shuthelah: of Eran, the

family of the Eranites.

rum 26:37 hae sunt cognationes filiorum Ephraim quarum 26:37 These are the families of the sons of Ephraim ac-

cording to those that were numbered of them, thirty and two thousand and five hundred. These are the sons of

numerus triginta duo milia quingentorum 26:38 isti sunt fi- Joseph after their families.

26:38 The sons of Benjamin after their families: of Bela, the family of the Belaites: of Ashbel, the family of the Ashbelites: of Ahiram, the family of the Ahiramites:

lii Ioseph per familias suas filii Beniamin in cognationibus suis Bale a quo familia Baleitarum Azbel a quo familia Azbelitarum Ahiram a quo familia Ahiramitarum 26:39 Su- 26:39 Of Shupham, the family of the Shuphamites: of

Hupham, the family of the Huphamites.

pham a quo familia Suphamitarum Hupham a quo familia Huphamitarum 26:40 filii Bale Hered et Noeman de Hered 26:40 And the sons of Bela were Ard and Naaman: of

Ard, the family of the Ardites: and of Naaman, the family of the Naamites.

familia Hereditarum de Noeman familia Noemitarum 26:41

26:41 These are the sons of Benjamin after their families: and they that were numbered of them were forty and five thousand and six hundred.

Biblia Sacra Vulgata

hii sunt filii Beniamin per cognationes suas quorum nume-

26:42 These are the sons of Dan after their families: of rus quadraginta quinque milia sescentorum 26:42 filii Dan

Shuham, the family of the Shuhamites. These are the families of Dan after their families.

per cognationes suas Suham a quo familia Suhamitarum hae

26:43 All the families of the Shuhamites, according to cognationes Dan per familias suas 26:43 omnes fuere Su-

those that were numbered of them, were threescore and four thousand and four hundred.

hamitae quorum numerus erat sexaginta quattuor milia qua-

26:44 Of the children of Asher after their families: of dringentorum 26:44 filii Aser per cognationes suas Iemna

Jimna, the family of the Jimnites: of Jesui, the family of the Jesuites: of Beriah, the family of the Beriites.

a quo familia Iemnaitarum Iessui a quo familia Iessuitarum

26:45 Of the sons of Beriah: of Heber, the family of the Brie a quo familia Brieitarum 26:45 filii Brie Haber a quo

Heberites: of Malchiel, the family of the Malchielites.

familia Haberitarum et Melchihel a quo familia Melchiheli-

26:46 And the name of the daughter of Asher was Sarah. tarum 26:46 nomen autem filiae Aser fuit Sara 26:47 hae

26:47 These are the families of the sons of Asher according to those that were numbered of them; who were fifty and three thousand and four hundred.

cognationes filiorum Aser et numerus eorum quinquaginta

26:48 Of the sons of Naphtali after their families: of tria milia quadringentorum 26:48 filii Nepthalim per cogna-

Jahzeel, the family of the Jahzeelites: of Guni, the family of the Gunites:

tiones suas Iessihel a quo familia Iessihelitarum Guni a quo

26:49 Of Jezer, the family of the Jezerites: of Shillem, the familia Gunitarum 26:49 Iesser a quo familia Iesseritarum family of the Shillemites. 26:50 These are the families of Naphtali according to their Sellem a quo familia Sellemitarum 26:50 hae sunt cognafamilies: and they that were numbered of them were forty and five thousand and four hundred.

tiones filiorum Nepthalim per familias suas quorum nume-

26:51 These were the numbered of the children of Israel, rus quadraginta quinque milia quadringentorum 26:51 ista

six hundred thousand and a thousand seven hundred and thirty.

est summa filiorum Israhel qui recensiti sunt sescenta milia

26:52 And the LORD spake unto Moses, saying, et mille septingenti triginta 26:52 locutusque est Dominus 26:53 Unto these the land shall be divided for an inheri- ad Mosen dicens 26:53 istis diuidetur terra iuxta numerum

tance according to the number of names.

26:54 To many thou shalt give the more inheritance, and uocabulorum in possessiones suas 26:54 pluribus maiorem

to few thou shalt give the less inheritance: to every one shall his inheritance be given according to those that were

numbered of him. partem dabis et paucioribus minorem singulis sicut nunc re-26:55 Notwithstanding the land shall be divided by lot: censiti sunt tradetur possessio 26:55 ita dumtaxat ut sors

according to the names of the tribes of their fathers they shall inherit.

terram tribubus diuidat et familiis 26:56 quicquid sorte con-

26:56 According to the lot shall the possession thereof be divided between many and few.

26:57 And these are they that were numbered of the tigerit hoc uel plures accipient uel pauciores 26:57 hic quo-

Levites after their families: of Gershon, the family of the Gershonites: of Kohath, the family of the Kohathites: of

Merari, the family of the Merarites. que est numerus filiorum Leui per familias suas Gerson a

quo familia Gersonitarum Caath a quo familia Caathitarum

26:58 These are the families of the Levites: the family of Merari a quo familia Meraritarum 26:58 hae sunt familiae the Libnites, the family of the Hebronites, the family of the Mahlites, the family of the Mushites, the family of the

Korathites. And Kohath begat Amram. Leui familia Lobni familia Hebroni familia Mooli familia 26:59 And the name of Amram’s wife was Jochebed, the Musi familia Cori at uero Caath genuit Amram 26:59 qui

daughter of Levi, whom her mother bare to Levi in Egypt: and she bare unto Amram Aaron and Moses, and Miriam

their sister. habuit uxorem Iochabed filiam Leui quae nata est ei in Ae-

gypto haec genuit uiro suo Amram filios Aaron et Mosen

26:60 And unto Aaron was born Nadab, and Abihu, et Mariam sororem eorum 26:60 de Aaron orti sunt Nadab

Eleazar, and Ithamar.

et Abiu et Eleazar et Ithamar 26:61 quorum Nadab et Abiu 26:61 And Nadab and Abihu died, when they offered

strange fire before the LORD.

mortui sunt cum obtulissent ignem alienum coram Domino 26:62 fueruntque omnes qui numerati sunt uiginti tria milia 26:62 And those that were numbered of them were twenty

and three thousand, all males from a month old and upward: for they were not numbered among the children of Israel, because there was no inheritance given them among the children of Israel.

generis masculini ab uno mense et supra quia non sunt recensiti inter filios Israhel nec eis cum ceteris data possessio 26:63 hic est numerus filiorum Israhel qui descripti sunt a 26:63 These are they that were numbered by Moses and

Eleazar the priest, who numbered the children of Israel in the plains of Moab by Jordan near Jericho.

Mosen et Eleazaro sacerdote in campestribus Moab supra Iordanem contra Hiericho 26:64 inter quos nullus fuit eo- 26:64 But among these there was not a man of them whom

Moses and Aaron the priest numbered, when they numbered the children of Israel in the wilderness of Sinai.

rum qui ante numerati sunt a Mose et Aaron in deserto Sinai 26:65 praedixerat enim Dominus quod omnes morerentur in 26:65 For the LORD had said of them, They shall surely

die in the wilderness. And there was not left a man of them, save Caleb the son of Jephunneh, and Joshua the

solitudine nullusque remansit ex eis nisi Chaleb filius Iep- son of Nun. phonne et Iosue filius Nun

27:1 accesserunt autem filiae Salphaad filii Epher filii Ga- 27:1 Then came the daughters of Zelophehad, the son of

Hepher, the son of Gilead, the son of Machir, the son of Manasseh, of the families of Manasseh the son of Joseph: and these are the names of his daughters; Mahlah, Noah, and Hoglah, and Milcah, and Tirzah.

laad filii Machir filii Manasse qui fuit filius Ioseph quarum

sunt nomina Maala et Noa et Egla et Melcha et Thersa 27:2 27:2 And they stood before Moses, and before Eleazar the priest, and before the princes and all the congregation, by the door of the tabernacle of the congregation, saying,

steteruntque coram Mosen et Eleazaro sacerdote et cunctis principibus populi ad ostium tabernaculi foederis atque dixerunt 27:3 pater noster mortuus est in deserto nec fuit in se- 27:3 Our father died in the wilderness, and he was not in

the company of them that gathered themselves together against the LORD in the company of Korah; but died in

ditione quae concitata est contra Dominum sub Core sed in his own sin, and had no sons. peccato suo mortuus est hic non habuit mares filios cur tollitur nomen illius de familia sua quia non habet filium date nobis possessionem inter cognatos patris nostri 27:4 rettu- 27:4 Why should the name of our father be done away

from among his family, because he hath no son? Give unto us therefore a possession among the brethren of our

litque Moses causam earum ad iudicium Domini 27:5 qui father.

27:5 And Moses brought their cause before the LORD.

dixit ad eum 27:6 iustam rem postulant filiae Salphaad da 27:6 And the LORD spake unto Moses, saying, eis possessionem inter cognatos patris sui et ei in hereditate succedant 27:7 ad filios autem Israhel loqueris haec 27:8 27:7 The daughters of Zelophehad speak right: thou shalt

surely give them a possession of an inheritance among their father’s brethren; and thou shalt cause the inheritance of their father to pass unto them. 27:8 And thou shalt speak unto the children of Israel, saying, If a man die, and have no son, then ye shall cause his inheritance to pass unto his daughter.

homo cum mortuus fuerit absque filio ad filiam eius transibit hereditas 27:9 si filiam non habuerit habebit successo-

27:9 And if he have no daughter, then ye shall give his

res fratres suos 27:10 quod si et fratres non fuerint dabitis _{inheritance unto his brethren.}

27:10 And if he have no brethren, then ye shall give his inheritance unto his father’s brethren.

hereditatem fratribus patris eius 27:11 sin autem nec pat-

27:11 And if his father have no brethren, then ye shall give his inheritance unto his kinsman that is next to him of his family, and he shall possess it: and it shall be unto the children of Israel a statute of judgment, as the LORD commanded Moses.

ruos habuerit dabitur hereditas his qui ei proximi sunt eritque hoc filiis Israhel sanctum lege perpetua sicut praecepit Dominus Mosi 27:12 dixit quoque Dominus ad Mosen as- 27:12 And the LORD said unto Moses, Get thee up into

this mount Abarim, and see the land which I have given unto the children of Israel.

Biblia Sacra Vulgata

cende in montem istum Abarim et contemplare inde terram

27:13 And when thou hast seen it, thou also shalt be gath- quam daturus sum filiis Israhel 27:13 cumque uideris eam ered unto thy people, as Aaron thy brother was gathered.

27:14 For ye rebelled against my commandment in the ibis et tu ad populum tuum sicut iuit frater tuus Aaron 27:14

desert of Zin, in the strife of the congregation, to sanctify me at the water before their eyes: that is the water of

Meribah in Kadesh in the wilderness of Zin. quia offendistis me in deserto Sin in contradictione multi-

tudinis nec sanctificare me uoluistis coram ea super aquas

27:15 And Moses spake unto the LORD, saying, hae sunt aquae Contradictionis in Cades deserti Sin 27:15 27:16 Let the LORD, the God of the spirits of all flesh, set cui respondit Moses 27:16 prouideat Dominus Deus spiri-

a man over the congregation,

tuum omnis carnis hominem qui sit super multitudinem hanc

27:17 Which may go out before them, and which may go 27:17 et possit exire et intrare ante eos et educere illos uel

in before them, and which may lead them out, and which may bring them in; that the congregation of the LORD be

not as sheep which have no shepherd. introducere ne sit populus Domini sicut oues absque pas-27:18 And the LORD said unto Moses, Take thee Joshua tore 27:18 dixitque Dominus ad eum tolle Iosue filium Nun

the son of Nun, a man in whom is the spirit, and lay thine hand upon him;

uirum in quo est spiritus et pone manum tuam super eum

27:19 And set him before Eleazar the priest, and before all 27:19 qui stabit coram Eleazaro sacerdote et omni multitu-

the congregation; and give him a charge in their sight.

27:20 And thou shalt put some of thine honour upon him, dine 27:20 et dabis ei praecepta cunctis uidentibus et par-

that all the congregation of the children of Israel may be obedient.

tem gloriae tuae ut audiat eum omnis synagoga filiorum Is-

27:21 And he shall stand before Eleazar the priest, who rahel 27:21 pro hoc si quid agendum erit Eleazar sacerdos

shall ask counsel for him after the judgment of Urim before the LORD: at his word shall they go out, and at his word they shall come in, both he, and all the children of Israel with him, even all the congregation.

consulet Dominum ad uerbum eius egredietur et ingredietur

27:22 And Moses did as the LORD commanded him: and ipse et omnes filii Israhel cum eo et cetera multitudo 27:22

he took Joshua, and set him before Eleazar the priest, and before all the congregation:

fecit Moses ut praeceperat Dominus cumque tulisset Iosue statuit eum coram Eleazaro sacerdote et omni frequentia po-

27:23 And he laid his hands upon him, and gave him a puli 27:23 et inpositis capiti eius manibus cuncta replicauit

charge, as the LORD commanded by the hand of Moses.

quae mandauerat Dominus

28:1 And the LORD spake unto Moses, saying, 28:1 dixit quoque Dominus ad Mosen 28:2 praecipe filiis

28:2 Command the children of Israel, and say unto them, My offering, and my bread for my sacrifices made by fire, for a sweet savour unto me, shall ye observe to offer unto me in their due season.

Israhel et dices ad eos oblationem meam et panes et incen-

28:3 And thou shalt say unto them, This is the offering sum odoris suauissimi offerte per tempora sua 28:3 haec

made by fire which ye shall offer unto the LORD; two lambs of the first year without spot day by day, for a con-

tinual burnt offering. sunt sacrificia quae offerre debetis agnos anniculos inmacu-28:4 The one lamb shalt thou offer in the morning, and the latos duos cotidie in holocaustum sempiternum 28:4 unum

other lamb shalt thou offer at even;

28:5 And a tenth part of an ephah of flour for a meat of- offeretis mane et alterum ad uesperam 28:5 decimam par-

fering, mingled with the fourth part of an hin of beaten oil.

tem oephi similae quae conspersa sit oleo purissimo et ha-

28:6 It is a continual burnt offering, which was ordained beat quartam partem hin 28:6 holocaustum iuge est quod in mount Sinai for a sweet savour, a sacrifice made by fire unto the LORD.

obtulistis in monte Sinai in odorem suauissimum incensi

28:7 And the drink offering thereof shall be the fourth part Domini 28:7 et libabitis uini quartam partem hin per agof an hin for the one lamb: in the holy place shalt thou cause the strong wine to be poured unto the LORD for a

drink offering. nos singulos in sanctuario Domini 28:8 alterumque agnum

28:8 And the other lamb shalt thou offer at even: as the meat offering of the morning, and as the drink offering thereof, thou shalt offer it, a sacrifice made by fire, of a sweet savour unto the LORD.

CCLXXII

similiter offeretis ad uesperam iuxta omnem ritum sacrificii matutini et libamentorum eius oblationem suauissimi odoris Domino 28:9 die autem sabbati offeretis duos agnos anni- 28:9 And on the sabbath day two lambs of the first year

without spot, and two tenth deals of flour for a meat offering, mingled with oil, and the drink offering thereof:

culos inmaculatos et duas decimas similae oleo conspersae in sacrificio et liba 28:10 quae rite funduntur per singula 28:10 This is the burnt offering of every sabbath, beside

the continual burnt offering, and his drink offering.

sabbata in holocausto sempiterno 28:11 in kalendis autem 28:11 And in the beginnings of your months ye shall offer

a burnt offering unto the LORD; two young bullocks, and one ram, seven lambs of the first year without spot;

id est in mensuum exordiis offeretis holocaustum Domino uitulos de armento duos arietem unum agnos anniculos septem inmaculatos 28:12 et tres decimas similae oleo con- 28:12 And three tenth deals of flour for a meat offering,

mingled with oil, for one bullock; and two tenth deals of flour for a meat offering, mingled with oil, for one ram;

spersae in sacrificio per singulos uitulos et duas decimas similae oleo conspersae per singulos arietes 28:13 et deci- 28:13 And a several tenth deal of flour mingled with oil

for a meat offering unto one lamb; for a burnt offering of a sweet savour, a sacrifice made by fire unto the LORD.

mam decimae similae ex oleo in sacrificio per agnos singulos holocaustum suauissimi odoris atque incensi est Domino 28:14 libamenta autem uini quae per singulas fundenda sunt 28:14 And their drink offerings shall be half an hin of wine

unto a bullock, and the third part of an hin unto a ram, and a fourth part of an hin unto a lamb: this is the burnt

uictimas ista erunt media pars hin per uitulos singulos tertia offering of every month throughout the months of the year. per arietem quarta per agnum hoc erit holocaustum per omnes menses qui sibi anno uertente succedunt 28:15 hircus 28:15 And one kid of the goats for a sin offering unto the

LORD shall be offered, beside the continual burnt offering, and his drink offering.

quoque offeretur Domino pro peccatis in holocaustum sempiternum cum libamentis suis 28:16 mense autem primo 28:16 And in the fourteenth day of the first month is the

passover of the LORD.

quartadecima die mensis phase Domini erit 28:17 et quin- 28:17 And in the fifteenth day of this month is the feast:

seven days shall unleavened bread be eaten.

tadecima die sollemnitas septem diebus uescentur azymis 28:18 quarum dies prima uenerabilis et sancta erit omne opus 28:18 In the first day shall be an holy convocation; ye shall

do no manner of servile work therein:

seruile non facietis in ea 28:19 offeretisque incensum holo- 28:19 But ye shall offer a sacrifice made by fire for a burnt

offering unto the LORD; two young bullocks, and one ram, and seven lambs of the first year: they shall be unto

caustum Domino uitulos de armento duos arietem unum ag- you without blemish: nos anniculos inmaculatos septem 28:20 et sacrificia singu- 28:20 And their meat offering shall be of flour mingled

with oil: three tenth deals shall ye offer for a bullock, and two tenth deals for a ram;

lorum ex simila quae conspersa sit oleo tres decimas per singulos uitulos et duas decimas per arietem 28:21 et decimam 28:21 A several tenth deal shalt thou offer for every lamb,

throughout the seven lambs:

decimae per agnos singulos id est per septem agnos 28:22 et 28:22 And one goat for a sin offering, to make an atone-

ment for you.

hircum pro peccato unum ut expietur pro uobis 28:23 prae- 28:23 Ye shall offer these beside the burnt offering in the

morning, which is for a continual burnt offering.

ter holocaustum matutinum quod semper offertis 28:24 ita 28:24 After this manner ye shall offer daily, throughout

the seven days, the meat of the sacrifice made by fire, of a sweet savour unto the LORD: it shall be offered beside

facietis per singulos dies septem dierum in fomitem ignis et the continual burnt offering, and his drink offering. in odorem suauissimum Domino qui surget de holocausto et de libationibus singulorum 28:25 dies quoque septimus 28:25 And on the seventh day ye shall have an holy con-

vocation; ye shall do no servile work.

Biblia Sacra Vulgata

celeberrimus et sanctus erit uobis omne opus seruile non fa-

28:26 Also in the day of the firstfruits, when ye bring a cietis in eo 28:26 dies etiam primitiuorum quando offertis

new meat offering unto the LORD, after your weeks be out, ye shall have an holy convocation; ye shall do no

servile work: nouas fruges Domino expletis ebdomadibus uenerabilis et 28:27 But ye shall offer the burnt offering for a sweet sancta erit omne opus seruile non facietis in ea 28:27 of-

savour unto the LORD; two young bullocks, one ram, seven lambs of the first year;

feretisque holocaustum in odorem suauissimum Domino uitulos de armento duos arietem unum et agnos anniculos in-

28:28 And their meat offering of flour mingled with oil, maculatos septem 28:28 atque in sacrificiis eorum similae

three tenth deals unto one bullock, two tenth deals unto one ram,

oleo conspersae tres decimas per singulos uitulos per arie-

28:29 A several tenth deal unto one lamb, throughout the tes duas 28:29 per agnos decimam decimae qui simul sunt seven lambs;

28:30 And one kid of the goats, to make an atonement for agni septem hircum quoque 28:30 qui mactatur pro expia-

you.

28:31 Ye shall offer them beside the continual burnt offer- tione praeter holocaustum sempiternum et liba eius 28:31

ing, and his meat offering, (they shall be unto you without blemish) and their drink offerings.

inmaculata offeretis omnia cum libationibus suis

29:1 And in the seventh month, on the first day of the 29:1 mensis etiam septimi prima dies uenerabilis et sancta

month, ye shall have an holy convocation; ye shall do no servile work: it is a day of blowing the trumpets unto you.

erit uobis omne opus seruile non facietis in ea quia dies clan-

29:2 And ye shall offer a burnt offering for a sweet savour goris est et tubarum 29:2 offeretisque holocaustum in odo-

unto the LORD; one young bullock, one ram, and seven lambs of the first year without blemish:

rem suauissimum Domino uitulum de armento unum arie-

29:3 And their meat offering shall be of flour mingled with tem unum agnos anniculos inmaculatos septem 29:3 et in oil, three tenth deals for a bullock, and two tenth deals for a ram,

sacrificiis eorum similae oleo conspersae tres decimas per

29:4 And one tenth deal for one lamb, throughout the singulos uitulos duas decimas per arietem 29:4 unam de-

seven lambs:

29:5 And one kid of the goats for a sin offering, to make cimam per agnum qui simul sunt agni septem 29:5 et hir-

an atonement for you:

29:6 Beside the burnt offering of the month, and his meat cum pro peccato qui offertur in expiationem populi 29:6

offering, and the daily burnt offering, and his meat offering, and their drink offerings, according unto their manner,

for a sweet savour, a sacrifice made by fire unto the LORD. praeter holocaustum kalendarum cum sacrificiis suis et holocaustum sempiternum cum libationibus solitis hisdem caerimoniis offeretis in odorem suauissimum incensum Domino

29:7 And ye shall have on the tenth day of this seventh 29:7 decima quoque dies mensis huius septimi erit uobis

month an holy convocation; and ye shall afflict your souls: ye shall not do any work therein:

sancta atque uenerabilis et adfligetis animas uestras omne

29:8 But ye shall offer a burnt offering unto the LORD for opus seruile non facietis in ea 29:8 offeretisque holocau-

a sweet savour; one young bullock, one ram, and seven lambs of the first year; they shall be unto you without

blemish: stum Domino in odorem suauissimum uitulum de armento

unum arietem unum agnos anniculos inmaculatos septem

29:9 And their meat offering shall be of flour mingled with 29:9 et in sacrificiis eorum similae oleo conspersae tres de-

oil, three tenth deals to a bullock, and two tenth deals to one ram,

cimas per uitulos singulos duas decimas per arietem 29:10

29:10 A several tenth deal for one lamb, throughout the seven lambs:

decimam decimae per agnos singulos qui sunt simul septem

29:11 One kid of the goats for a sin offering; beside the agni 29:11 et hircum pro peccato absque his quae offerri

sin offering of atonement, and the continual burnt offering, and the meat offering of it, and their drink offerings.

pro delicto solent in expiationem et holocaustum sempiternum in sacrificio et libaminibus eorum 29:12 quintadecima 29:12 And on the fifteenth day of the seventh month ye

shall have an holy convocation; ye shall do no servile work, and ye shall keep a feast unto the LORD seven days:

uero die mensis septimi quae uobis erit sancta atque uenerabilis omne opus seruile non facietis in ea sed celebrabitis sollemnitatem Domino septem diebus 29:13 offeretisque 29:13 And ye shall offer a burnt offering, a sacrifice made

by fire, of a sweet savour unto the LORD; thirteen young bullocks, two rams, and fourteen lambs of the first year;

holocaustum in odorem suauissimum Domino uitulos de ar- they shall be without blemish: mento tredecim arietes duos agnos anniculos quattuordecim inmaculatos 29:14 et in libamentis eorum similae oleo con- 29:14 And their meat offering shall be of flour mingled

with oil, three tenth deals unto every bullock of the thirteen bullocks, two tenth deals to each ram of the two rams,

spersae tres decimas per uitulos singulos qui sunt simul uituli tredecim et duas decimas arieti uno id est simul arieti-

bus duobus 29:15 et decimam decimae agnis singulis qui 29:15 And a several tenth deal to each lamb of the fourteen

lambs:

sunt simul agni quattuordecim 29:16 et hircum pro peccato 29:16 And one kid of the goats for a sin offering; be-

side the continual burnt offering, his meat offering, and his drink offering.

absque holocausto sempiterno et sacrificio et libamine eius 29:17 in die altero offeres uitulos de armento duodecim arie- 29:17 And on the second day ye shall offer twelve young

bullocks, two rams, fourteen lambs of the first year without spot:

tes duos agnos anniculos inmaculatos quattuordecim 29:18

29:18 And their meat offering and their drink offerings for the bullocks, for the rams, and for the lambs, shall be according to their number, after the manner:

sacrificiaque et libamina singulorum per uitulos et arietes et agnos rite celebrabis 29:19 et hircum pro peccato absque 29:19 And one kid of the goats for a sin offering; beside

the continual burnt offering, and the meat offering thereof, and their drink offerings.

holocausto sempiterno sacrificioque eius et libamine 29:20

29:20 And on the third day eleven bullocks, two rams, fourteen lambs of the first year without blemish;

die tertio offeres uitulos undecim arietes duos agnos anniculos inmaculatos quattuordecim 29:21 sacrificiaque et li- 29:21 And their meat offering and their drink offerings

for the bullocks, for the rams, and for the lambs, shall be according to their number, after the manner:

bamina singulorum per uitulos et arietes et agnos rite celebrabis 29:22 et hircum pro peccato absque holocausto sem- 29:22 And one goat for a sin offering; beside the continual

burnt offering, and his meat offering, and his drink offering.

piterno et sacrificio et libamine eius 29:23 die quarto offe-

29:23 And on the fourth day ten bullocks, two rams, and fourteen lambs of the first year without blemish:

res uitulos decem arietes duos agnos anniculos inmaculatos quattuordecim 29:24 sacrificiaque eorum et libamina sin- 29:24 Their meat offering and their drink offerings for the

bullocks, for the rams, and for the lambs, shall be according to their number, after the manner:

gulorum per uitulos et arietes et agnos rite celebrabis 29:25

29:25 And one kid of the goats for a sin offering; beside the continual burnt offering, his meat offering, and his drink offering.

et hircum pro peccato absque holocausto sempiterno sacrificioque eius et libamine 29:26 die quinto offeres uitulos no- 29:26 And on the fifth day nine bullocks, two rams, and

fourteen lambs of the first year without spot:

uem arietes duos agnos anniculos inmaculatos quattuordecim 29:27 sacrificiaque et libamina singulorum per uitulos 29:27 And their meat offering and their drink offerings

for the bullocks, for the rams, and for the lambs, shall be according to their number, after the manner:

et arietes et agnos rite celebrabis 29:28 et hircum pro pec-

29:28 And one goat for a sin offering; beside the continual burnt offering, and his meat offering, and his drink offering.

cato absque holocausto sempiterno sacrificioque eius et libamine 29:29 die sexto offeres uitulos octo arietes duos agnos 29:29 And on the sixth day eight bullocks, two rams, and

fourteen lambs of the first year without blemish:

Biblia Sacra Vulgata

29:30 And their meat offering and their drink offerings anniculos inmaculatos quattuordecim 29:30 sacrificiaque et

for the bullocks, for the rams, and for the lambs, shall be according to their number, after the manner:

libamina singulorum per uitulos et arietes et agnos rite cele-

29:31 And one goat for a sin offering; beside the continual brabis 29:31 et hircum pro peccato absque holocausto sem-

burnt offering, his meat offering, and his drink offering.

29:32 And on the seventh day seven bullocks, two rams, piterno sacrificioque eius et libamine 29:32 die septimo of-

and fourteen lambs of the first year without blemish:

feres uitulos septem arietes duos agnos anniculos inmacu-

29:33 And their meat offering and their drink offerings latos quattuordecim 29:33 sacrificiaque et libamina singu-

for the bullocks, for the rams, and for the lambs, shall be according to their number, after the manner:

lorum per uitulos et arietes et agnos rite celebrabis 29:34 et

29:34 And one goat for a sin offering; beside the continual burnt offering, his meat offering, and his drink offering.

hircum pro peccato absque holocausto sempiterno sacrificio-

29:35 On the eighth day ye shall have a solemn assembly: que eius et libamine 29:35 die octauo qui est celeberrimus

ye shall do no servile work therein:

29:36 But ye shall offer a burnt offering, a sacrifice made omne opus seruile non facietis 29:36 offerentes holocaus-

by fire, of a sweet savour unto the LORD: one bullock, one ram, seven lambs of the first year without blemish:

tum in odorem suauissimum Domino uitulum unum arietem

29:37 Their meat offering and their drink offerings for the unum agnos anniculos inmaculatos septem 29:37 sacrificia-

bullock, for the ram, and for the lambs, shall be according to their number, after the manner:

que et libamina singulorum per uitulos et arietes et agnos rite

29:38 And one goat for a sin offering; beside the continual celebrabis 29:38 et hircum pro peccato absque holocausto

burnt offering, and his meat offering, and his drink offering.

sempiterno sacrificioque eius et libamine 29:39 haec offe-

29:39 These things ye shall do unto the LORD in your set feasts, beside your vows, and your freewill offerings, for your burnt offerings, and for your meat offerings, and for your drink offerings, and for your peace offerings.

retis Domino in sollemnitatibus uestris praeter uota et oblationes spontaneas in holocausto in sacrificio in libamine et in hostiis pacificis

30:1 And Moses told the children of Israel according to all 30:1 narrauitque Moses filiis Israhel omnia quae ei Domithat the LORD commanded Moses.

30:2 And Moses spake unto the heads of the tribes con- nus imperarat 30:2 et locutus est ad principes tribuum fi-

cerning the children of Israel, saying, This is the thing which the LORD hath commanded.

liorum Israhel iste est sermo quem praecepit Dominus 30:3

30:3 If a man vow a vow unto the LORD, or swear an oath to bind his soul with a bond; he shall not break his word, he shall do according to all that proceedeth out of his mouth.

si quis uirorum uotum Domino uouerit aut se constrinxerit iuramento non faciet irritum uerbum suum sed omne quod

30:4 If a woman also vow a vow unto the LORD, and bind promisit implebit 30:4 mulier si quippiam uouerit et se con-

herself by a bond, being in her father’s house in her youth;

strinxerit iuramento quae est in domo patris sui et in aetate adhuc puellari si cognouerit pater uotum quod pollicita est et iuramentum quo obligauit animam suam et tacuerit uoti

30:5 And her father hear her vow, and her bond where- rea erit 30:5 quicquid pollicita est et iurauit opere conplebit

with she hath bound her soul, and her father shall hold his peace at her; then all her vows shall stand, and every bond

wherewith she hath bound her soul shall stand. 30:6 sin autem statim ut audierit contradixerit pater et uota

30:6 But if her father disallow her in the day that he heareth; not any of her vows, or of her bonds wherewith she hath bound her soul, shall stand: and the LORD shall forgive her, because her father disallowed her.

et iuramenta eius irrita erunt nec obnoxia tenebitur spon-

30:7 And if she had at all an husband, when she vowed, sioni eo quod contradixerit pater 30:7 si maritum habue-

or uttered ought out of her lips, wherewith she bound her soul;

rit et uouerit aliquid et semel uerbum de ore eius egrediens

30:8 And her husband heard it, and held his peace at her animam illius obligauerit iuramento 30:8 quo die audierit

in the day that he heard it: then her vows shall stand, and her bonds wherewith she bound her soul shall stand.

uir et non contradixerit uoti rea erit reddet quodcumque promiserat 30:9 sin autem audiens statim contradixerit et irri- 30:9 But if her husband disallowed her on the day that he

heard it; then he shall make her vow which she vowed, and that which she uttered with her lips, wherewith she bound

tas fecerit pollicitationes eius uerbaque quibus obstrinxerat her soul, of none effect: and the LORD shall forgive her. animam suam propitius ei erit Dominus 30:10 uidua et re- 30:10 But every vow of a widow, and of her that is di-

vorced, wherewith they have bound their souls, shall stand against her.

pudiata quicquid uouerint reddent 30:11 uxor in domo uiri

30:11 And if she vowed in her husband’s house, or bound her soul by a bond with an oath;

cum se uoto constrinxerit et iuramento 30:12 si audierit uir _{30:12 And her husband heard it, and held his peace at her,}

and disallowed her not: then all her vows shall stand, and every bond wherewith she bound her soul shall stand.

et tacuerit nec contradixerit sponsioni reddet quodcumque promiserat 30:13 sin autem extemplo contradixerit non te- 30:13 But if her husband hath utterly made them void on

the day he heard them; then whatsoever proceeded out of her lips concerning her vows, or concerning the bond of her soul, shall not stand: her husband hath made them void; and the LORD shall forgive her.

nebitur promissionis rea quia maritus contradixit et Dominus ei propitius erit 30:14 si uouerit et iuramento se constrinxe- 30:14 Every vow, and every binding oath to afflict the soul,

her husband may establish it, or her husband may make it void.

rit ut per ieiunium uel ceterarum rerum abstinentiam adfligat animam suam in arbitrio uiri erit ut faciat siue non faciat 30:15 quod si audiens uir tacuerit et in alteram diem distule- 30:15 But if her husband altogether hold his peace at her

from day to day; then he establisheth all her vows, or all her bonds, which are upon her: he confirmeth them, because he held his peace at her in the day that he heard them.

rit sententiam quicquid uouerat atque promiserat reddet quia statim ut audiuit tacuit 30:16 sin autem contradixerit post- 30:16 But if he shall any ways make them void after that

he hath heard them; then he shall bear her iniquity.

quam resciuit portabit ipse iniquitatem eius 30:17 istae sunt 30:17 These are the statutes, which the LORD com-

manded Moses, between a man and his wife, between the father and his daughter, being yet in her youth in her fa-

leges quas constituit Dominus Mosi inter uirum et uxorem ther’s house. inter patrem et filiam quae in puellari adhuc aetate est uel quae manet in parentis domo

31:1 locutusque est Dominus ad Mosen dicens 31:2 ul- 31:1 And the LORD spake unto Moses, saying,

31:2 Avenge the children of Israel of the Midianites: afterward shalt thou be gathered unto thy people.

ciscere prius filios Israhel de Madianitis et sic colligeris ad populum tuum 31:3 statimque Moses armate inquit ex uo- 31:3 And Moses spake unto the people, saying, Arm some

of yourselves unto the war, and let them go against the Midianites, and avenge the LORD of Midian.

bis uiros ad pugnam qui possint ultionem Domini expetere de Madianitis 31:4 mille uiri de singulis tribubus eligan- 31:4 Of every tribe a thousand, throughout all the tribes of

Israel, shall ye send to the war.

tur Israhel qui mittantur ad bellum 31:5 dederuntque mille- 31:5 So there were delivered out of the thousands of Israel,

a thousand of every tribe, twelve thousand armed for war.

nos de cunctis tribubus id est duodecim milia expeditorum

ad pugnam 31:6 quos misit Moses cum Finees filio Elea- 31:6 And Moses sent them to the war, a thousand of every tribe, them and Phinehas the son of Eleazar the priest, to the war, with the holy instruments, and the trumpets to

zari sacerdotis uasa quoque sancta et tubas ad clangendum blow in his hand. tradidit ei 31:7 cumque pugnassent contra Madianitas at- 31:7 And they warred against the Midianites, as the LORD

commanded Moses; and they slew all the males.

que uicissent omnes mares occiderunt 31:8 et reges eorum 31:8 And they slew the kings of Midian, beside the rest of

them that were slain; namely, Evi, and Rekem, and Zur, and Hur, and Reba, five kings of Midian: Balaam also the

Eui et Recem et Sur et Vr et Rebe quinque principes gentis son of Beor they slew with the sword. Balaam quoque filium Beor interfecerunt gladio 31:9 ce- 31:9 And the children of Israel took all the women of Mid-

ian captives, and their little ones, and took the spoil of all their cattle, and all their flocks, and all their goods.

peruntque mulieres eorum et paruulos omniaque pecora et

Biblia Sacra Vulgata

cunctam supellectilem quicquid habere potuerant depopulati

31:10 And they burnt all their cities wherein they dwelt, sunt 31:10 tam urbes quam uiculos et castella flamma conand all their goodly castles, with fire. 31:11 And they took all the spoil, and all the prey, both of sumpsit 31:11 et tulerunt praedam et uniuersa quae cepemen and of beasts.

31:12 And they brought the captives, and the prey, and rant tam ex hominibus quam ex iumentis 31:12 et adduxe-

the spoil, unto Moses, and Eleazar the priest, and unto the congregation of the children of Israel, unto the camp at the

plains of Moab, which are by Jordan near Jericho. runt ad Mosen et Eleazarum sacerdotem et ad omnem multitudinem filiorum Israhel reliqua etiam utensilia portauerunt ad castra in campestribus Moab iuxta Iordanem contra Hie-

31:13 And Moses, and Eleazar the priest, and all the richo 31:13 egressi sunt autem Moses et Eleazar sacerdos

princes of the congregation, went forth to meet them without the camp.

et omnes principes synagogae in occursum eorum extra cas-

31:14 And Moses was wroth with the officers of the host, tra 31:14 iratusque Moses principibus exercitus tribunis et

with the captains over thousands, and captains over hundreds, which came from the battle.

centurionibus qui uenerant de bello 31:15 ait cur feminas

31:15 And Moses said unto them, Have ye saved all the women alive?

31:16 Behold, these caused the children of Israel, through reseruastis 31:16 nonne istae sunt quae deceperunt filios Is-

the counsel of Balaam, to commit trespass against the LORD in the matter of Peor, and there was a plague among

the congregation of the LORD. rahel ad suggestionem Balaam et praeuaricari uos fecerunt in

Domino super peccato Phogor unde et percussus est populus

31:17 Now therefore kill every male among the little ones, 31:17 ergo cunctos interficite quicquid est generis masculini

and kill every woman that hath known man by lying with him.

etiam in paruulis et mulieres quae nouerunt uiros in coitu iu-

31:18 But all the women children, that have not known a gulate 31:18 puellas autem et omnes feminas uirgines reser-

man by lying with him, keep alive for yourselves.

31:19 And do ye abide without the camp seven days: uate uobis 31:19 et manete extra castra septem diebus qui

whosoever hath killed any person, and whosoever hath touched any slain, purify both yourselves and your cap-

tives on the third day, and on the seventh day. occiderit hominem uel occisum tetigerit lustrabitur die tertio

31:20 And purify all your raiment, and all that is made of et septimo 31:20 et de omni praeda siue uestimentum fueskins, and all work of goats’ hair, and all things made of wood.

rit siue uas et aliquid in utensilia praeparatum de caprarum

31:21 And Eleazar the priest said unto the men of war pellibus et pilis et ligno expiabitur 31:21 Eleazar quoque sa-

which went to the battle, This is the ordinance of the law which the LORD commanded Moses;

cerdos ad uiros exercitus qui pugnauerant sic locutus est hoc

31:22 Only the gold, and the silver, the brass, the iron, the est praeceptum legis quod mandauit Dominus Mosi 31:22 tin, and the lead,

aurum et argentum et aes et ferrum et stagnum et plumbum

31:23 Every thing that may abide the fire, ye shall make 31:23 et omne quod potest transire per flammas igne purga-

it go through the fire, and it shall be clean: nevertheless it shall be purified with the water of separation: and all that

abideth not the fire ye shall make go through the water. bitur quicquid autem ignem non potest sustinere aqua ex-31:24 And ye shall wash your clothes on the seventh day, piationis sanctificabitur 31:24 et lauabitis uestimenta uestra

and ye shall be clean, and afterward ye shall come into the camp.

die septimo et purificati postea castra intrabitis 31:25 dixit-

31:25 And the LORD spake unto Moses, saying,

31:26 Take the sum of the prey that was taken, both of que Dominus ad Mosen 31:26 tollite summam eorum quae

man and of beast, thou, and Eleazar the priest, and the chief fathers of the congregation:

capta sunt ab homine usque ad pecus tu et Eleazar sacerdos

31:27 And divide the prey into two parts; between them et principes uulgi 31:27 diuidesque ex aequo praedam inter

that took the war upon them, who went out to battle, and between all the congregation:

eos qui pugnauerunt et egressi sunt ad bellum et inter omnem reliquam multitudinem 31:28 et separabis partem Domino 31:28 And levy a tribute unto the Lord of the men of war

which went out to battle: one soul of five hundred, both of the persons, and of the beeves, and of the asses, and of the

ab his qui pugnauerunt et fuerunt in bello unam animam de sheep: quingentis tam ex hominibus quam ex bubus et asinis et ouibus 31:29 et dabis eam Eleazaro sacerdoti quia primitiae 31:29 Take it of their half, and give it unto Eleazar the

priest, for an heave offering of the LORD.

Domini sunt 31:30 ex media quoque parte filiorum Israhel 31:30 And of the children of Israel’s half, thou shalt take one portion of fifty, of the persons, of the beeves, of the asses, and of the flocks, of all manner of beasts, and give them unto the Levites, which keep the charge of the tabernacle of the LORD.

accipies quinquagesimum caput hominum et boum et asinorum et ouium cunctarumque animantium et dabis ea Leuitis qui excubant in custodiis tabernaculi Domini 31:31 fece- 31:31 And Moses and Eleazar the priest did as the LORD

commanded Moses.

runtque Moses et Eleazar sicut praeceperat Dominus 31:32 31:32 And the booty, being the rest of the prey which the

men of war had caught, was six hundred thousand and seventy thousand and five thousand sheep,

fuit autem praeda quam exercitus ceperat ouium sescenta septuaginta quinque milia 31:33 boum septuaginta duo milia 31:33 And threescore and twelve thousand beeves, 31:34 asinorum sexaginta milia et mille 31:35 animae homi- 31:34 And threescore and one thousand asses,

31:35 And thirty and two thousand persons in all, of women that had not known man by lying with him.

num sexus feminei quae non cognouerant uiros triginta duo milia 31:36 dataque est media pars his qui in proelio fue- 31:36 And the half, which was the portion of them that

went out to war, was in number three hundred thousand and seven and thirty thousand and five hundred sheep:

rant ouium trecenta triginta septem milia quingenta 31:37

31:37 And the LORD’s tribute of the sheep was six hundred and threescore and fifteen.

e quibus in partem Domini supputatae sunt oues sescentae septuaginta quinque 31:38 et de bubus triginta sex milibus 31:38 And the beeves were thirty and six thousand; of

which the LORD’s tribute was threescore and twelve.

boues septuaginta duo 31:39 de asinis triginta milibus quin- 31:39 And the asses were thirty thousand and five hun-

dred; of which the LORD’s tribute was threescore and one.

gentis asini sexaginta unus 31:40 de animabus hominum 31:40 And the persons were sixteen thousand; of which

the LORD’s tribute was thirty and two persons.

sedecim milibus cesserunt in partem Domini triginta duae animae 31:41 tradiditque Moses numerum primitiarum Do- 31:41 And Moses gave the tribute, which was the LORD’s

heave offering, unto Eleazar the priest, as the LORD commanded Moses.

mini Eleazaro sacerdoti sicut ei fuerat imperatum 31:42 ex

31:42 And of the children of Israel’s half, which Moses divided from the men that warred,

media parte filiorum Israhel quam separauerat his qui in proelio fuerant 31:43 de media uero parte quae contigerat re- 31:43 (Now the half that pertained unto the congregation

was three hundred thousand and thirty thousand and seven thousand and five hundred sheep,

liquae multitudini id est de ouium trecentis triginta septem milibus quingentis 31:44 et de bubus triginta sex milibus 31:44 And thirty and six thousand beeves, 31:45 et de asinis triginta milibus quingentis 31:46 et de 31:45 And thirty thousand asses and five hundred,

31:46 And sixteen thousand persons;)

hominibus sedecim milibus 31:47 tulit Moses quinquagesi- _{31:47 Even of the children of Israel’s half, Moses took one}

portion of fifty, both of man and of beast, and gave them unto the Levites, which kept the charge of the tabernacle

mum caput et dedit Leuitis qui excubant in tabernaculo Do- _{of the LORD; as the LORD commanded Moses.} mini sicut praeceperat Dominus 31:48 cumque accessissent 31:48 And the officers which were over thousands of the

host, the captains of thousands, and captains of hundreds, came near unto Moses:

principes exercitus ad Mosen et tribuni centurionesque dixerunt 31:49 nos serui tui recensuimus numerum pugnatorum 31:49 And they said unto Moses, Thy servants have taken

the sum of the men of war which are under our charge, and there lacketh not one man of us.

quos habuimus sub manu nostra et ne unus quidem defuit

Biblia Sacra Vulgata

31:50 We have therefore brought an oblation for the 31:50 ob hanc causam offerimus in donariis Domini singuli

LORD, what every man hath gotten, of jewels of gold, chains, and bracelets, rings, earrings, and tablets, to make

an atonement for our souls before the LORD. quod in praeda auri potuimus inuenire periscelides et armil-

las anulos et dextralia ac murenulas ut depreceris pro nobis

31:51 And Moses and Eleazar the priest took the gold of Dominum 31:51 susceperuntque Moses et Eleazar sacerdos

them, even all wrought jewels.

31:52 And all the gold of the offering that they offered omne aurum in diuersis speciebus 31:52 pondo sedecim mi-

up to the LORD, of the captains of thousands, and of the captains of hundreds, was sixteen thousand seven hundred

and fifty shekels. lia septingentos quinquaginta siclos a tribunis et centurioni-31:53 (For the men of war had taken spoil, every man for bus 31:53 unusquisque enim quod in praeda rapuerat suum

himself.)

31:54 And Moses and Eleazar the priest took the gold of erat 31:54 et susceptum intulerunt in tabernaculum testimo-

the captains of thousands and of hundreds, and brought it into the tabernacle of the congregation, for a memorial for

the children of Israel before the LORD. nii in monumentum filiorum Israhel coram Domino

32:1 Now the children of Reuben and the children of Gad 32:1 filii autem Ruben et Gad habebant pecora multa et had a very great multitude of cattle: and when they saw the land of Jazer, and the land of Gilead, that, behold, the

place was a place for cattle; erat illis in iumentis infinita substantia cumque uidissent Ia-32:2 The children of Gad and the children of Reuben came zer et Galaad aptas alendis animalibus 32:2 uenerunt ad

and spake unto Moses, and to Eleazar the priest, and unto the princes of the congregation, saying,

Mosen et ad Eleazarum sacerdotem et principes multitudi-

32:3 Ataroth, and Dibon, and Jazer, and Nimrah, and Hes- nis atque dixerunt 32:3 Atharoth et Dibon et Iazer et Nemra

hbon, and Elealeh, and Shebam, and Nebo, and Beon, 32:4 Even the country which the LORD smote before the Esbon et Eleale et Sabam et Nebo et Beon 32:4 terram congregation of Israel, is a land for cattle, and thy servants have cattle:

quam percussit Dominus in conspectu filiorum Israhel regionis uberrimae est ad pastum animalium et nos serui tui

32:5 Wherefore, said they, if we have found grace in thy habemus iumenta plurima 32:5 precamurque si inuenimus

sight, let this land be given unto thy servants for a possession, and bring us not over Jordan.

gratiam coram te ut des nobis famulis tuis eam in possessio-

32:6 And Moses said unto the children of Gad and to the nem ne facias nos transire Iordanem 32:6 quibus respon-

children of Reuben, Shall your brethren go to war, and shall ye sit here?

dit Moses numquid fratres uestri ibunt ad pugnam et uos

32:7 And wherefore discourage ye the heart of the children hic sedebitis 32:7 cur subuertitis mentes filiorum Israhel

of Israel from going over into the land which the LORD hath given them?

ne transire audeant in locum quem eis daturus est Dominus

32:8 Thus did your fathers, when I sent them from 32:8 nonne ita egerunt patres uestri quando misi de Cades-

Kadeshbarnea to see the land. 32:9 For when they went up unto the valley of Eshcol, and barne ad explorandam terram 32:9 cumque uenissent ussaw the land, they discouraged the heart of the children of Israel, that they should not go into the land which the

LORD had given them. que ad uallem Botri lustrata omni regione subuerterunt cor

filiorum Israhel ut non intrarent fines quos eis Dominus de-

32:10 And the LORD’s anger was kindled the same time, dit 32:10 qui iratus iurauit dicens 32:11 si uidebunt homi-

and he sware, saying, 32:11 Surely none of the men that came up out of Egypt, from twenty years old and upward, shall see the land which I sware unto Abraham, unto Isaac, and unto Jacob; because they have not wholly followed me:

nes isti qui ascenderunt ex Aegypto a uiginti annis et supra terram quam sub iuramento pollicitus sum Abraham Isaac

32:12 Save Caleb the son of Jephunneh the Kenezite, and et Iacob et noluerunt sequi me 32:12 praeter Chaleb filium

Joshua the son of Nun: for they have wholly followed the LORD.

Iepphonne Cenezeum et Iosue filium Nun isti impleuerunt

32:13 And the LORD’s anger was kindled against Israel, uoluntatem meam 32:13 iratusque Dominus aduersum Isra-

and he made them wander in the wilderness forty years, until all the generation, that had done evil in the sight of the LORD, was consumed.

hel circumduxit eum per desertum quadraginta annis donec consumeretur uniuersa generatio quae fecerat malum in conspectu eius 32:14 et ecce inquit uos surrexistis pro patribus 32:14 And, behold, ye are risen up in your fathers’ stead,

an increase of sinful men, to augment yet the fierce anger of the LORD toward Israel.

uestris incrementa et alumni hominum peccatorum ut augeretis furorem Domini contra Israhel 32:15 qui si nolueritis 32:15 For if ye turn away from after him, he will yet again

leave them in the wilderness; and ye shall destroy all this people.

sequi eum in solitudine populum derelinquet et uos causa eritis necis omnium 32:16 at illi prope accedentes dixerunt 32:16 And they came near unto him, and said, We will

build sheepfolds here for our cattle, and cities for our little ones:

caulas ouium fabricabimus et stabula iumentorum paruulis quoque nostris urbes munitas 32:17 nos autem ipsi armati 32:17 But we ourselves will go ready armed before the

children of Israel, until we have brought them unto their place: and our little ones shall dwell in the fenced cities

et accincti pergemus ad proelium ante filios Israhel donec because of the inhabitants of the land. introducamus eos ad loca sua paruuli nostri et quicquid habere possumus erunt in urbibus muratis propter habitatorum insidias 32:18 non reuertemur in domos nostras usquequo 32:18 We will not return unto our houses, until the chil-

dren of Israel have inherited every man his inheritance.

possideant filii Israhel hereditatem suam 32:19 nec quic- 32:19 For we will not inherit with them on yonder side

Jordan, or forward; because our inheritance is fallen to us on this side Jordan eastward.

quam quaeremus trans Iordanem quia iam habemus possessionem nostram in orientali eius plaga 32:20 quibus Moses 32:20 And Moses said unto them, If ye will do this thing,

if ye will go armed before the LORD to war,

ait si facitis quod promittitis expediti pergite coram Domino ad pugnam 32:21 et omnis uir bellator armatus Iordanem 32:21 And will go all of you armed over Jordan before the

LORD, until he hath driven out his enemies from before him,

transeat donec subuertat Dominus inimicos suos 32:22 et

32:22 And the land be subdued before the LORD: then afterward ye shall return, and be guiltless before the LORD, and before Israel; and this land shall be your possession before the LORD.

subiciatur ei omnis terra tunc eritis inculpabiles et apud Dominum et apud Israhel et obtinebitis regiones quas uultis coram Domino 32:23 sin autem quod dicitis non feceritis nulli 32:23 But if ye will not do so, behold, ye have sinned

against the LORD: and be sure your sin will find you out.

dubium quin peccetis in Dominum et scitote quoniam peccatum uestrum adprehendet uos 32:24 aedificate ergo urbes 32:24 Build you cities for your little ones, and folds for

your sheep; and do that which hath proceeded out of your mouth.

paruulis uestris et caulas ac stabula ouibus ac iumentis et quod polliciti estis implete 32:25 dixeruntque filii Gad et 32:25 And the children of Gad and the children of Reuben

spake unto Moses, saying, Thy servants will do as my lord commandeth.

Ruben ad Mosen serui tui sumus faciemus quod iubet dominus noster 32:26 paruulos nostros et mulieres et pecora ac 32:26 Our little ones, our wives, our flocks, and all our

cattle, shall be there in the cities of Gilead:

iumenta relinquemus in urbibus Galaad 32:27 nos autem fa- 32:27 But thy servants will pass over, every man armed

for war, before the LORD to battle, as my lord saith.

muli tui omnes expediti pergemus ad bellum sicut tu domine loqueris 32:28 praecepit ergo Moses Eleazaro sacerdoti et 32:28 So concerning them Moses commanded Eleazar the

priest, and Joshua the son of Nun, and the chief fathers of the tribes of the children of Israel:

Iosue filio Nun et principibus familiarum per tribus Israhel et dixit ad eos 32:29 si transierint filii Gad et filii Ruben 32:29 And Moses said unto them, If the children of Gad

and the children of Reuben will pass with you over Jordan, every man armed to battle, before the LORD, and the land shall be subdued before you; then ye shall give them the land of Gilead for a possession:

Biblia Sacra Vulgata

uobiscum Iordanem omnes armati ad bellum coram Domino et uobis fuerit terra subiecta date eis Galaad in possessio-

32:30 But if they will not pass over with you armed, they nem 32:30 sin autem noluerint transire uobiscum in terram

shall have possessions among you in the land of Canaan.

32:31 And the children of Gad and the children of Reuben Chanaan inter uos habitandi accipiant loca 32:31 respon-

answered, saying, As the LORD hath said unto thy servants, so will we do.

deruntque filii Gad et filii Ruben sicut locutus est Dominus

32:32 We will pass over armed before the LORD into the seruis suis ita faciemus 32:32 ipsi armati pergemus coram

land of Canaan, that the possession of our inheritance on this side Jordan may be ours.

Domino in terram Chanaan et possessionem iam suscepisse

32:33 And Moses gave unto them, even to the children of nos confitemur trans Iordanem 32:33 dedit itaque Moses fi-

Gad, and to the children of Reuben, and unto half the tribe of Manasseh the son of Joseph, the kingdom of Sihon king of the Amorites, and the kingdom of Og king of Bashan, the land, with the cities thereof in the coasts, even the cities of the country round about.

liis Gad et Ruben et dimidiae tribui Manasse filii Ioseph regnum Seon regis Amorrei et regnum Og regis Basan et terram

32:34 And the children of Gad built Dibon, and Ataroth, eorum cum urbibus suis per circuitum 32:34 igitur extruxe-

and Aroer,

32:35 And Atroth, Shophan, and Jaazer, and Jogbehah, runt filii Gad Dibon et Atharoth et Aroer 32:35 Etrothso-

32:36 And Bethnimrah, and Bethharan, fenced cities: and phan et Iazer Iecbaa 32:36 et Bethnemra et Betharan urbes folds for sheep.

32:37 And the children of Reuben built Heshbon, and munitas et caulas pecoribus suis 32:37 filii uero Ruben ae-

Elealeh, and Kirjathaim,

32:38 And Nebo, and Baalmeon, (their names being dificauerunt Esbon et Eleale et Cariathaim 32:38 et Nabo

changed,) and Shibmah: and gave other names unto the cities which they builded.

et Baalmeon uersis nominibus Sabama quoque inponentes

32:39 And the children of Machir the son of Manasseh uocabula urbibus quas extruxerant 32:39 porro filii Machir

went to Gilead, and took it, and dispossessed the Amorite which was in it.

filii Manasse perrexerunt in Galaad et uastauerunt eam inter-

32:40 And Moses gave Gilead unto Machir the son of fecto Amorreo habitatore eius 32:40 dedit ergo Moses ter-

Manasseh; and he dwelt therein.

32:41 And Jair the son of Manasseh went and took the ram Galaad Machir filio Manasse qui habitauit in ea 32:41

small towns thereof, and called them Havothjair.

Iair autem filius Manasse abiit et occupauit uicos eius quos

32:42 And Nobah went and took Kenath, and the villages appellauit Auothiair id est uillas Iair 32:42 Nobe quoque

thereof, and called it Nobah, after his own name.

perrexit et adprehendit Canath cum uiculis suis uocauitque eam ex nomine suo Nobe

33:1 These are the journeys of the children of Israel, which 33:1 hae sunt mansiones filiorum Israhel qui egressi sunt went forth out of the land of Egypt with their armies under the hand of Moses and Aaron.

de Aegypto per turmas suas in manu Mosi et Aaron 33:2

33:2 And Moses wrote their goings out according to their journeys by the commandment of the LORD: and these are their journeys according to their goings out.

quas descripsit Moses iuxta castrorum loca quae Domini

33:3 And they departed from Rameses in the first month, iussione mutabant 33:3 profecti igitur de Ramesse mense

on the fifteenth day of the first month; on the morrow after the passover the children of Israel went out with an high

hand in the sight of all the Egyptians. primo quintadecima die mensis primi altera die phase filii Is-

33:4 For the Egyptians buried all their firstborn, which the rahel in manu excelsa uidentibus cunctis Aegyptiis 33:4 et LORD had smitten among them: upon their gods also the LORD executed judgments.

sepelientibus primogenitos quos percusserat Dominus nam

33:5 And the children of Israel removed from Rameses, et in diis eorum exercuerat ultionem 33:5 castrametati sunt

and pitched in Succoth.

33:6 And they departed from Succoth, and pitched in in Soccoth 33:6 et de Soccoth uenerunt in Aetham quae est

Etham, which is in the edge of the wilderness.

in extremis finibus solitudinis 33:7 inde egressi uenerunt 33:7 And they removed from Etham, and turned again

unto Pihahiroth, which is before Baalzephon: and they pitched before Migdol.

contra Phiahiroth quae respicit Beelsephon et castrametati sunt ante Magdolum 33:8 profectique de Phiahiroth trans- 33:8 And they departed from before Pihahiroth, and

passed through the midst of the sea into the wilderness, and went three days’ journey in the wilderness of Etham,

ierunt per medium mare in solitudinem et ambulantes tribus and pitched in Marah. diebus per desertum Aetham castrametati sunt in Mara 33:9 33:9 And they removed from Marah, and came unto Elim:

and in Elim were twelve fountains of water, and threescore and ten palm trees; and they pitched there.

profectique de Mara uenerunt in Helim ubi erant duodecim fontes aquarum et palmae septuaginta ibique castrametati sunt 33:10 sed et inde egressi fixere tentoria super mare Ru- 33:10 And they removed from Elim, and encamped by the

Red sea.

brum profectique de mari Rubro 33:11 castrametati sunt in 33:11 And they removed from the Red sea, and encamped

in the wilderness of Sin.

deserto Sin 33:12 unde egressi uenerunt in Dephca 33:13 33:12 And they took their journey out of the wilderness of Sin, and encamped in Dophkah. 33:13 And they departed from Dophkah, and encamped in Alush.

profectique de Dephca castrametati sunt in Alus 33:14 eg-

33:14 And they removed from Alush, and encamped at

ressi de Alus Raphidim fixere tentoria ubi aqua populo defuit ^{Rephidim, where was no water for the people to drink.} ad bibendum 33:15 profectique de Raphidim castrametati 33:15 And they departed from Rephidim, and pitched in

the wilderness of Sinai.

sunt in deserto Sinai 33:16 sed et de solitudine Sinai egressi 33:16 And they removed from the desert of Sinai, and

pitched at Kibrothhattaavah.

uenerunt ad sepulchra Concupiscentiae 33:17 profectique 33:17 And they departed from Kibrothhattaavah, and en-

camped at Hazeroth.

de sepulchris Concupiscentiae castrametati sunt in Aseroth 33:18 et de Aseroth uenerunt in Rethma 33:19 profectique 33:18 And they departed from Hazeroth, and pitched in

Rithmah.

33:19 And they departed from Rithmah, and pitched at Rimmonparez.

de Rethma castrametati sunt in Remmonphares 33:20 unde

33:20 And they departed from Rimmonparez, and pitched

egressi uenerunt in Lebna 33:21 et de Lebna castrametati ^{in Libnah.}

33:21 And they removed from Libnah, and pitched at Rissah.

sunt in Ressa 33:22 egressi de Ressa uenerunt in Ceelatha

33:22 And they journeyed from Rissah, and pitched in Kehelathah.

33:23 unde profecti castrametati sunt in monte Sepher 33:24 _{33:23 And they went from Kehelathah, and pitched in}

mount Shapher.

33:24 And they removed from mount Shapher, and encamped in Haradah.

egressi de monte Sepher uenerunt in Arada 33:25 inde pro-

33:25 And they removed from Haradah, and pitched in Makheloth.

ficiscentes castrametati sunt in Maceloth 33:26 profectique

33:26 And they removed from Makheloth, and encamped

de Maceloth uenerunt in Thaath 33:27 de Thaath castra- ^{at Tahath.}

33:27 And they departed from Tahath, and pitched at Tarah.

metati sunt in Thare 33:28 unde egressi fixerunt tentoria

33:28 And they removed from Tarah, and pitched in Mithcah.

in Methca 33:29 et de Methca castrametati sunt in Esmona _{33:29 And they went from Mithcah, and pitched in Hash-}

monah.

33:30 profectique de Esmona uenerunt in Moseroth 33:31 et 33:30 And they departed from Hashmonah, and encamped

at Moseroth.

33:31 And they departed from Moseroth, and pitched in Benejaakan.

de Moseroth castrametati sunt in Baneiacan 33:32 egressi-

33:32 And they removed from Benejaakan, and encamped

que de Baneiacan uenerunt in montem Gadgad 33:33 unde ^{at Horhagidgad.}

33:33 And they went from Horhagidgad, and pitched in Jotbathah.

profecti castrametati sunt in Hietebatha 33:34 et de Hiete-

33:34 And they removed from Jotbathah, and encamped at Ebronah.

batha uenerunt in Ebrona 33:35 egressique de Ebrona ca-

33:35 And they departed from Ebronah, and encamped at Eziongaber.

strametati sunt in Asiongaber 33:36 inde profecti uenerunt 33:36 And they removed from Eziongaber, and pitched in

the wilderness of Zin, which is Kadesh.

in desertum Sin haec est Cades 33:37 egressique de Ca- 33:37 And they removed from Kadesh, and pitched in

mount Hor, in the edge of the land of Edom.

des castrametati sunt in monte Hor in extremis finibus ter-

33:38 And Aaron the priest went up into mount Hor at the rae Edom 33:38 ascenditque Aaron sacerdos montem Hor commandment of the LORD, and died there, in the fortieth year after the children of Israel were come out of the land

of Egypt, in the first day of the fifth month. iubente Domino et ibi mortuus est anno quadragesimo egressionis filiorum Israhel ex Aegypto mense quinto prima

33:39 And Aaron was an hundred and twenty and three die mensis 33:39 cum esset annorum centum uiginti trium

years old when he died in mount Hor.

33:40 And king Arad the Canaanite, which dwelt in the 33:40 audiuitque Chananeus rex Arad qui habitabat ad me-

south in the land of Canaan, heard of the coming of the children of Israel.

ridiem in terra Chanaan uenisse filios Israhel 33:41 et pro-

33:41 And they departed from mount Hor, and pitched in Zalmonah.

33:42 And they departed from Zalmonah, and pitched in fecti de monte Hor castrametati sunt in Salmona 33:42 unde

Punon.

33:43 And they departed from Punon, and pitched in egressi uenerunt in Phinon 33:43 profectique de Phinon ca-

Oboth.

33:44 And they departed from Oboth, and pitched in Ije- strametati sunt in Oboth 33:44 et de Oboth uenerunt in Ie-

abarim, in the border of Moab.

33:45 And they departed from Iim, and pitched in Dibon- abarim quae est in finibus Moabitarum 33:45 profectique

gad.

33:46 And they removed from Dibongad, and encamped de Ieabarim fixere tentoria in Dibongad 33:46 unde egressi

in Almondiblathaim.

33:47 And they removed from Almondiblathaim, and castrametati sunt in Elmondeblathaim 33:47 egressi de El-

pitched in the mountains of Abarim, before Nebo.

mondeblathaim uenerunt ad montes Abarim contra Nabo

33:48 And they departed from the mountains of Abarim, 33:48 profectique de montibus Abarim transierunt ad cam-

and pitched in the plains of Moab by Jordan near Jericho.

33:49 And they pitched by Jordan, from Bethjesimoth pestria Moab super Iordanem contra Hiericho 33:49 ibique

even unto Abelshittim in the plains of Moab.

castrametati sunt de Bethsimon usque ad Belsattim in pla-

33:50 And the LORD spake unto Moses in the plains of nioribus locis Moabitarum 33:50 ubi locutus est Dominus

Moab by Jordan near Jericho, saying, 33:51 Speak unto the children of Israel, and say unto them, ad Mosen 33:51 praecipe filiis Israhel et dic ad eos quando When ye are passed over Jordan into the land of Canaan;

33:52 Then ye shall drive out all the inhabitants of the land transieritis Iordanem intrantes terram Chanaan 33:52 dis-

from before you, and destroy all their pictures, and destroy all their molten images, and quite pluck down all their high

places: perdite cunctos habitatores regionis illius confringite titulos

33:53 And ye shall dispossess the inhabitants of the land, et statuas comminuite atque omnia excelsa uastate 33:53 and dwell therein: for I have given you the land to possess it.

mundantes terram et habitantes in ea ego enim dedi uobis

33:54 And ye shall divide the land by lot for an inheritance illam in possessionem 33:54 quam diuidetis uobis sorte plu-

among your families: and to the more ye shall give the more inheritance, and to the fewer ye shall give the less inheritance: every man’s inheritance shall be in the place where his lot falleth; according to the tribes of your fathers ye shall inherit.

ribus dabitis latiorem et paucis angustiorem singulis ut sors ceciderit ita tribuetur hereditas per tribus et familias posses-

33:55 But if ye will not drive out the inhabitants of the sio diuidetur 33:55 sin autem nolueritis interficere habita-

land from before you; then it shall come to pass, that those which ye let remain of them shall be pricks in your eyes, and thorns in your sides, and shall vex you in the land wherein ye dwell.

tores terrae qui remanserint erunt uobis quasi claui in oculis et lanceae in lateribus et aduersabuntur uobis in terra habita-

33:56 Moreover it shall come to pass, that I shall do unto tionis uestrae 33:56 et quicquid illis facere cogitaram uobis

you, as I thought to do unto them.

faciam

34:1 And the LORD spake unto Moses, saying, 34:1 locutus est Dominus ad Mosen 34:2 praecipe filiis

34:2 Command the children of Israel, and say unto them, When ye come into the land of Canaan; (this is the land that shall fall unto you for an inheritance, even the land of Canaan with the coasts thereof:)

Israhel et dices ad eos cum ingressi fueritis terram Chanaan et in possessionem uobis sorte ceciderit his finibus terminabitur 34:3 pars meridiana incipiet a solitudine Sin quae est 34:3 Then your south quarter shall be from the wilderness

of Zin along by the coast of Edom, and your south border shall be the outmost coast of the salt sea eastward:

iuxta Edom et habebit terminos contra orientem mare Salsissimum 34:4 qui circumibunt australem plagam per as- 34:4 And your border shall turn from the south to the as-

cent of Akrabbim, and pass on to Zin: and the going forth thereof shall be from the south to Kadeshbarnea, and shall

censum Scorpionis ita ut transeant Senna et perueniant in go on to Hazaraddar, and pass on to Azmon: meridiem usque ad Cadesbarne unde egredientur confinia ad uillam nomine Addar et tendent usque Asemona 34:5 34:5 And the border shall fetch a compass from Azmon

unto the river of Egypt, and the goings out of it shall be at the sea.

ibitque per gyrum terminus ab Asemona usque ad torrentem Aegypti et maris Magni litore finietur 34:6 plaga autem oc- 34:6 And as for the western border, ye shall even have the

great sea for a border: this shall be your west border.

cidentalis a mari Magno incipiet et ipso fine cludetur 34:7 34:7 And this shall be your north border: from the great

sea ye shall point out for you mount Hor:

porro ad septentrionalem plagam a mari Magno termini incipient peruenientes usque ad montem Altissimum 34:8 a 34:8 From mount Hor ye shall point out your border unto

the entrance of Hamath; and the goings forth of the border shall be to Zedad:

quo uenies in Emath usque ad terminos Sedada 34:9 ibunt-

34:9 And the border shall go on to Ziphron, and the goings out of it shall be at Hazarenan: this shall be your north border.

que confinia usque Zephrona et uillam Henan hii erunt termini in parte aquilonis 34:10 inde metabuntur fines contra 34:10 And ye shall point out your east border from Haz-

arenan to Shepham:

orientalem plagam de uilla Henan usque Sephama 34:11 et 34:11 And the coast shall go down from Shepham to Ri-

blah, on the east side of Ain; and the border shall descend, and shall reach unto the side of the sea of Chinnereth east-

de Sephama descendent termini in Rebla contra fontem in- ward: de peruenient contra orientem ad mare Chenereth 34:12 et 34:12 And the border shall go down to Jordan, and the

goings out of it shall be at the salt sea: this shall be your land with the coasts thereof round about.

tendent usque Iordanem et ad ultimum Salsissimo cludentur mari hanc habebitis terram per fines suos in circuitu 34:13 34:13 And Moses commanded the children of Israel, say-

ing, This is the land which ye shall inherit by lot, which the LORD commanded to give unto the nine tribes, and to

praecepitque Moses filiis Israhel dicens haec erit terra quam the half tribe: possidebitis sorte et quam iussit dari Dominus nouem tribubus et dimidiae tribui 34:14 tribus enim filiorum Ruben 34:14 For the tribe of the children of Reuben according to

the house of their fathers, and the tribe of the children of Gad according to the house of their fathers, have received their inheritance; and half the tribe of Manasseh have received their inheritance:

per familias suas et tribus filiorum Gad iuxta cognationum

numerum media quoque tribus Manasse 34:15 id est duae 34:15 The two tribes and the half tribe have received their inheritance on this side Jordan near Jericho eastward, toward the sunrising.

semis tribus acceperunt partem suam trans Iordanem contra Hiericho ad orientalem plagam 34:16 et ait Dominus ad 34:16 And the LORD spake unto Moses, saying, Mosen 34:17 haec sunt nomina uirorum qui terram uobis 34:17 These are the names of the men which shall divide

the land unto you: Eleazar the priest, and Joshua the son of Nun.

diuident Eleazar sacerdos et Iosue filius Nun 34:18 et sin-

34:18 And ye shall take one prince of every tribe, to divide the land by inheritance.

guli principes de tribubus singulis 34:19 quorum ista sunt 34:19 And the names of the men are these: Of the tribe of

Judah, Caleb the son of Jephunneh.

uocabula de tribu Iuda Chaleb filius Iepphonne 34:20 de 34:20 And of the tribe of the children of Simeon, Shemuel

the son of Ammihud.

tribu Symeon Samuhel filius Ammiud 34:21 de tribu Be- 34:21 Of the tribe of Benjamin, Elidad the son of Chislon.

niamin Helidad filius Chaselon 34:22 de tribu filiorum Dan 34:22 And the prince of the tribe of the children of Dan,

Bukki the son of Jogli.

Biblia Sacra Vulgata

34:23 The prince of the children of Joseph, for the tribe of Bocci filius Iogli 34:23 filiorum Ioseph de tribu Manasse the children of Manasseh, Hanniel the son of Ephod.

34:24 And the prince of the tribe of the children of Hannihel filius Ephod 34:24 de tribu Ephraim Camuhel fi-

Ephraim, Kemuel the son of Shiphtan.

34:25 And the prince of the tribe of the children of Zebu- lius Sephtan 34:25 de tribu Zabulon Elisaphan filius Phar-

lun, Elizaphan the son of Parnach.

34:26 And the prince of the tribe of the children of Is- nach 34:26 de tribu Isachar dux Faltihel filius Ozan 34:27

sachar, Paltiel the son of Azzan. 34:27 And the prince of the tribe of the children of Asher, Ahihud the son of Shelomi.

de tribu Aser Ahiud filius Salomi 34:28 de tribu Nepthali

34:28 And the prince of the tribe of the children of Naph-

tali, Pedahel the son of Ammihud. Phedahel filius Ameiud 34:29 hii sunt quibus praecepit Do-

34:29 These are they whom the LORD commanded to divide the inheritance unto the children of Israel in the land of Canaan.

minus ut diuiderent filiis Israhel terram Chanaan

35:1 And the LORD spake unto Moses in the plains of 35:1 haec quoque locutus est Dominus ad Mosen in cam-

Moab by Jordan near Jericho, saying,

35:2 Command the children of Israel, that they give unto pestribus Moab super Iordanem contra Hiericho 35:2 prae-

the Levites of the inheritance of their possession cities to dwell in; and ye shall give also unto the Levites suburbs

for the cities round about them. cipe filiis Israhel ut dent Leuitis de possessionibus suis 35:3

35:3 And the cities shall they have to dwell in; and the suburbs of them shall be for their cattle, and for their goods, and for all their beasts.

urbes ad habitandum et suburbana earum per circuitum ut ipsi in oppidis maneant et suburbana sint pecoribus ac iu-

35:4 And the suburbs of the cities, which ye shall give mentis 35:4 quae a muris ciuitatum forinsecus per circui-

unto the Levites, shall reach from the wall of the city and outward a thousand cubits round about.

tum mille passuum spatio tendentur 35:5 contra orientem

35:5 And ye shall measure from without the city on the east side two thousand cubits, and on the south side two thousand cubits, and on the west side two thousand cubits, and on the north side two thousand cubits; and the city shall be in the midst: this shall be to them the suburbs of the cities.

duo milia erunt cubiti et contra meridiem similiter duo milia ad mare quoque quod respicit occidentem eadem mensura erit et septentrionalis plaga aequali termino finietur erunt-

35:6 And among the cities which ye shall give unto the que urbes in medio et foris suburbana 35:6 de ipsis autem

Levites there shall be six cities for refuge, which ye shall appoint for the manslayer, that he may flee thither: and to

them ye shall add forty and two cities. oppidis quae Leuitis dabitis sex erunt in fugitiuorum auxi-

lia separata ut fugiat ad ea qui fuderit sanguinem exceptis

35:7 So all the cities which ye shall give to the Levites his alia quadraginta duo oppida 35:7 id est simul quadra-

shall be forty and eight cities: them shall ye give with their suburbs.

ginta octo cum suburbanis suis 35:8 ipsaeque urbes quae

35:8 And the cities which ye shall give shall be of the possession of the children of Israel: from them that have many ye shall give many; but from them that have few ye shall give few: every one shall give of his cities unto the Levites according to his inheritance which he inheriteth.

dabuntur de possessionibus filiorum Israhel ab his qui plus habent plures auferentur et qui minus pauciores singuli iuxta

35:9 And the LORD spake unto Moses, saying, mensuram hereditatis suae dabunt oppida Leuitis 35:9 ait

35:10 Speak unto the children of Israel, and say unto them, Dominus ad Mosen 35:10 loquere filiis Israhel et dices ad When ye be come over Jordan into the land of Canaan;

eos quando transgressi fueritis Iordanem in terram Chanaan

35:11 Then ye shall appoint you cities to be cities of refuge 35:11 decernite quae urbes esse debeant in praesidia fugitiuo-

for you; that the slayer may flee thither, which killeth any person at unawares.

rum qui nolentes sanguinem fuderint 35:12 in quibus cum

35:12 And they shall be unto you cities for refuge from the avenger; that the manslayer die not, until he stand before the congregation in judgment.

fuerit profugus cognatus occisi eum non poterit occidere donec stet in conspectu multitudinis et causa illius iudicetur

35:13 And of these cities which ye shall give six cities 35:13 de ipsis autem urbibus quae ad fugitiuorum subsidia

shall ye have for refuge.

35:14 Ye shall give three cities on this side Jordan, and separantur 35:14 tres erunt trans Iordanem et tres in terra

three cities shall ye give in the land of Canaan, which shall be cities of refuge.

Chanaan 35:15 tam filiis Israhel quam aduenis atque pereg- 35:15 These six cities shall be a refuge, both for the chil-

dren of Israel, and for the stranger, and for the sojourner among them: that every one that killeth any person un-

rinis ut confugiat ad eas qui nolens sanguinem fuderit 35:16 awares may flee thither.

35:16 And if he smite him with an instrument of iron, so that he die, he is a murderer: the murderer shall surely be put to death.

si quis ferro percusserit et mortuus fuerit qui percussus est reus erit homicidii et ipse morietur 35:17 si lapidem iecerit 35:17 And if he smite him with throwing a stone, where-

with he may die, and he die, he is a murderer: the murderer shall surely be put to death.

et ictus occubuerit similiter punietur 35:18 si ligno percus-

35:18 Or if he smite him with an hand weapon of wood, wherewith he may die, and he die, he is a murderer: the murderer shall surely be put to death.

sus interierit percussoris sanguine uindicabitur 35:19 pro-

35:19 The revenger of blood himself shall slay the murderer: when he meeteth him, he shall slay him.

pinquus occisi homicidam interficiet statim ut adprehenderit eum percutiet 35:20 si per odium quis hominem inpule- 35:20 But if he thrust him of hatred, or hurl at him by

laying of wait, that he die;

rit uel iecerit quippiam in eum per insidias 35:21 aut cum 35:21 Or in enmity smite him with his hand, that he die:

he that smote him shall surely be put to death; for he is a murderer: the revenger of blood shall slay the murderer,

esset inimicus manu percusserit et ille mortuus fuerit per- when he meeteth him. cussor homicidii reus erit cognatus occisi statim ut inuenerit eum iugulabit 35:22 quod si fortuito et absque odio 35:23 35:22 But if he thrust him suddenly without enmity, or

have cast upon him any thing without laying of wait, 35:23 Or with any stone, wherewith a man may die, seeing him not, and cast it upon him, that he die, and was not his enemy, neither sought his harm:

et inimicitiis quicquam horum fecerit 35:24 et hoc audiente

35:24 Then the congregation shall judge between the slayer and the revenger of blood according to these judgments:

populo fuerit conprobatum atque inter percussorem et propinquum sanguinis quaestio uentilata 35:25 liberabitur in-

35:25 And the congregation shall deliver the slayer out of the hand of the revenger of blood, and the congregation shall restore him to the city of his refuge, whither he was fled: and he shall abide in it unto the death of the high priest, which was anointed with the holy oil.

nocens de ultoris manu et reducetur per sententiam in urbem ad quam confugerat manebitque ibi donec sacerdos magnus qui oleo sancto unctus est moriatur 35:26 si interfector ex- 35:26 But if the slayer shall at any time come without the

border of the city of his refuge, whither he was fled;

tra fines urbium quae exulibus deputatae sunt 35:27 fuerit 35:27 And the revenger of blood find him without the bor-

ders of the city of his refuge, and the revenger of blood kill the slayer; he shall not be guilty of blood:

inuentus et percussus ab eo qui ultor est sanguinis absque noxa erit qui eum occiderit 35:28 debuerat enim profugus 35:28 Because he should have remained in the city of his

refuge until the death of the high priest: but after the death of the high priest the slayer shall return into the land of his

usque ad mortem pontificis in urbe residere postquam autem possession. ille obierit homicida reuertetur in terram suam 35:29 haec 35:29 So these things shall be for a statute of judg-

ment unto you throughout your generations in all your dwellings.

sempiterna erunt et legitima in cunctis habitationibus uestris 35:30 homicida sub testibus punietur ad unius testimonium 35:30 Whoso killeth any person, the murderer shall be put

to death by the mouth of witnesses: but one witness shall not testify against any person to cause him to die.

nullus condemnabitur 35:31 non accipietis pretium ab eo

35:31 Moreover ye shall take no satisfaction for the life of a murderer, which is guilty of death: but he shall be surely put to death.

qui reus est sanguinis statim et ipse morietur 35:32 exules

35:32 And ye shall take no satisfaction for him that is fled to the city of his refuge, that he should come again to dwell

et profugi ante mortem pontificis nullo modo in urbes suas _{in the land, until the death of the priest.} reuerti poterunt 35:33 ne polluatis terram habitationis ue- 35:33 So ye shall not pollute the land wherein ye are: for

blood it defileth the land: and the land cannot be cleansed of the blood that is shed therein, but by the blood of him

strae quae insontium cruore maculatur nec aliter expiari po- that shed it. test nisi per eius sanguinem qui alterius sanguinem fuderit 35:34 atque ita emundabitur uestra possessio me commorante 35:34 Defile not therefore the land which ye shall inhabit,

wherein I dwell: for I the LORD dwell among the children of Israel.

uobiscum ego enim sum Dominus qui habito inter filios Is-

rahel

36:1 And the chief fathers of the families of the children 36:1 accesserunt autem et principes familiarum Galaad filii

of Gilead, the son of Machir, the son of Manasseh, of the families of the sons of Joseph, came near, and spake before Moses, and before the princes, the chief fathers of the children of Israel:

Machir filii Manasse de stirpe filiorum Ioseph locutique sunt

36:2 And they said, The LORD commanded my lord to Mosi coram principibus Israhel atque dixerunt 36:2 tibi do-

give the land for an inheritance by lot to the children of Israel: and my lord was commanded by the LORD to give the inheritance of Zelophehad our brother unto his daughters.

mino nostro praecepit Dominus ut terram sorte diuideres filiis Israhel et ut filiabus Salphaad fratris nostri dares pos-

36:3 And if they be married to any of the sons of the other sessionem debitam patri 36:3 quas si alterius tribus homi-

tribes of the children of Israel, then shall their inheritance be taken from the inheritance of our fathers, and shall be put to the inheritance of the tribe whereunto they are received: so shall it be taken from the lot of our inheritance.

nes uxores acceperint sequetur possessio sua et translata ad

36:4 And when the jubile of the children of Israel shall be, aliam tribum de nostra hereditate minuetur 36:4 atque ita then shall their inheritance be put unto the inheritance of the tribe whereunto they are received: so shall their inheritance be taken away from the inheritance of the tribe of our fathers.

fiet ut cum iobeleus id est quinquagesimus annus remissionis aduenerit confundatur sortium distributio et aliorum pos-

36:5 And Moses commanded the children of Israel accord- sessio ad alios transeat 36:5 respondit Moses filiis Israhel

ing to the word of the LORD, saying, The tribe of the sons of Joseph hath said well.

et Domino praecipiente ait recte tribus filiorum Ioseph lo-

36:6 This is the thing which the LORD doth command cuta est 36:6 et haec lex super filiabus Salphaad a Domino

concerning the daughters of Zelophehad, saying, Let them marry to whom they think best; only to the family of the

tribe of their father shall they marry. promulgata est nubant quibus uolunt tantum ut suae tribus 36:7 So shall not the inheritance of the children of Israel hominibus 36:7 ne commisceatur possessio filiorum Israhel

remove from tribe to tribe: for every one of the children of Israel shall keep himself to the inheritance of the tribe of

his fathers. de tribu in tribum omnes enim uiri ducent uxores de tribu et

36:8 And every daughter, that possesseth an inheritance in cognatione sua 36:8 et cunctae feminae maritos de eadem any tribe of the children of Israel, shall be wife unto one of the family of the tribe of her father, that the children of

Israel may enjoy every man the inheritance of his fathers. tribu accipient ut hereditas permaneat in familiis 36:9 nec

36:9 Neither shall the inheritance remove from one tribe to another tribe; but every one of the tribes of the children of Israel shall keep himself to his own inheritance.

sibi misceantur tribus sed ita maneant 36:10 ut a Domino

36:10 Even as the LORD commanded Moses, so did the

daughters of Zelophehad: separatae sunt feceruntque filiae Salphaad ut fuerat impera-36:11 For Mahlah, Tirzah, and Hoglah, and Milcah, and tum 36:11 et nupserunt Maala et Thersa et Egla et Melcha

Noah, the daughters of Zelophehad, were married unto their father’s brothers’ sons:

et Noa filiis patrui sui 36:12 de familia Manasse qui fuit fi-

36:12 And they were married into the families of the sons of Manasseh the son of Joseph, and their inheritance remained in the tribe of the family of their father.

lius Ioseph et possessio quae illis fuerat adtributa mansit in

36:13 These are the commandments and the judgments, tribu et familia patris earum 36:13 haec sunt mandata atque

which the LORD commanded by the hand of Moses unto the children of Israel in the plains of Moab by Jordan near

Jericho. iudicia quae praecepit Dominus per manum Mosi ad filios Israhel in campestribus Moab super Iordanem contra Hiericho

Liber Deuteronomii

1:1 haec sunt uerba quae locutus est Moses ad omnem Is- 1:1 These be the words which Moses spake unto all Is-

rael on this side Jordan in the wilderness, in the plain over against the Red sea, between Paran, and Tophel, and La-

rahel trans Iordanem in solitudine campestri contra mare Ru- ban, and Hazeroth, and Dizahab. brum inter Pharan et Thophel et Laban et Aseroth ubi auri est plurimum 1:2 undecim diebus de Horeb per uiam montis 1:2 (There are eleven days’ journey from Horeb by the

way of mount Seir unto Kadeshbarnea.)

Seir usque Cadesbarne 1:3 quadragesimo anno undecimo 1:3 And it came to pass in the fortieth year, in the eleventh month, on the first day of the month, that Moses spake unto the children of Israel, according unto all that the

mense prima die mensis locutus est Moses ad filios Israhel LORD had given him in commandment unto them; omnia quae praeceperat illi Dominus ut diceret eis 1:4 post- 1:4 After he had slain Sihon the king of the Amorites,

which dwelt in Heshbon, and Og the king of Bashan, which dwelt at Astaroth in Edrei:

quam percussit Seon regem Amorreorum qui habitauit in Esebon et Og regem Basan qui mansit in Aseroth et in Edrai 1:5 trans Iordanem in terra Moab coepitque Moses explanare 1:5 On this side Jordan, in the land of Moab, began Moses

to declare this law, saying,

legem et dicere 1:6 Dominus Deus noster locutus est ad 1:6 The LORD our God spake unto us in Horeb, saying,

Ye have dwelt long enough in this mount:

nos in Horeb dicens sufficit uobis quod in hoc monte mansistis 1:7 reuertimini et uenite ad montem Amorreorum et 1:7 Turn you, and take your journey, and go to the mount

of the Amorites, and unto all the places nigh thereunto, in the plain, in the hills, and in the vale, and in the south, and by the sea side, to the land of the Canaanites, and unto Lebanon, unto the great river, the river Euphrates.

ad cetera quae ei proxima sunt campestria atque montana et humiliora loca contra meridiem et iuxta litus maris terram Chananeorum et Libani usque ad flumen magnum Eufraten 1:8 en inquit tradidi uobis ingredimini et possidete eam su- 1:8 Behold, I have set the land before you: go in and pos-

sess the land which the LORD sware unto your fathers, Abraham, Isaac, and Jacob, to give unto them and to their

per qua iurauit Dominus patribus uestris Abraham et Isaac seed after them. et Iacob ut daret illam eis et semini eorum post eos 1:9 1:9 And I spake unto you at that time, saying, I am not

able to bear you myself alone:

dixique uobis illo in tempore 1:10 non possum solus su- 1:10 The LORD your God hath multiplied you, and, be-

hold, ye are this day as the stars of heaven for multitude.

stinere uos quia Dominus Deus uester multiplicauit uos et estis hodie sicut stellae caeli plurimae 1:11 Dominus De- 1:11 (The LORD God of your fathers make you a thousand

times so many more as ye are, and bless you, as he hath promised you!)

us patrum uestrorum addat ad hunc numerum multa milia et benedicat uobis sicut locutus est 1:12 non ualeo solus ues- 1:12 How can I myself alone bear your cumbrance, and

your burden, and your strife?

tra negotia sustinere et pondus ac iurgia 1:13 date e uobis 1:13 Take you wise men, and understanding, and known

among your tribes, and I will make them rulers over you.

uiros sapientes et gnaros et quorum conuersatio sit probata in tribubus uestris ut ponam eos uobis principes 1:14 tunc 1:14 And ye answered me, and said, The thing which thou

hast spoken is good for us to do.

respondistis mihi bona res est quam uis facere 1:15 tuli- 1:15 So I took the chief of your tribes, wise men, and

known, and made them heads over you, captains over thousands, and captains over hundreds, and captains over fifties, and captains over tens, and officers among your tribes.

que de tribubus uestris uiros sapientes et nobiles et constitui eos principes tribunos et centuriones et quinquagenarios ac decanos qui docerent uos singula 1:16 praecepique eis di- 1:16 And I charged your judges at that time, saying, Hear

the causes between your brethren, and judge righteously between every man and his brother, and the stranger that

cens audite illos et quod iustum est iudicate siue ciuis sit ille is with him. siue peregrinus 1:17 nulla erit distantia personarum ita par- 1:17 Ye shall not respect persons in judgment; but ye shall

hear the small as well as the great; ye shall not be afraid of the face of man; for the judgment is God’s: and the cause that is too hard for you, bring it unto me, and I will hear it.

Biblia Sacra Vulgata

uum audietis ut magnum nec accipietis cuiusquam personam quia Dei iudicium est quod si difficile uobis aliquid uisum

1:18 And I commanded you at that time all the things fuerit referte ad me et ego audiam 1:18 praecepique omnia

which ye should do.

1:19 And when we departed from Horeb, we went through quae facere deberetis 1:19 profecti autem de Horeb trans-

all that great and terrible wilderness, which ye saw by the way of the mountain of the Amorites, as the LORD our

God commanded us; and we came to Kadeshbarnea. iuimus per heremum terribilem et maximam quam uidistis

per uiam montis Amorrei sicut praeceperat Dominus Deus

1:20 And I said unto you, Ye are come unto the mountain noster nobis cumque uenissemus in Cadesbarne 1:20 dixi

of the Amorites, which the LORD our God doth give unto us.

uobis uenistis ad montem Amorrei quem Dominus Deus no-

1:21 Behold, the LORD thy God hath set the land before ster daturus est nobis 1:21 uide terram quam Dominus Deus

thee: go up and possess it, as the LORD God of thy fathers hath said unto thee; fear not, neither be discouraged.

tuus dat tibi ascende et posside eam sicut locutus est Dominus Deus patribus tuis noli metuere nec quicquam paueas

1:22 And ye came near unto me every one of you, and 1:22 et accessistis ad me omnes atque dixistis mittamus ui-

said, We will send men before us, and they shall search us out the land, and bring us word again by what way we

must go up, and into what cities we shall come. ros qui considerent terram et renuntient per quod iter debe-1:23 And the saying pleased me well: and I took twelve amus ascendere et ad quas pergere ciuitates 1:23 cumque

men of you, one of a tribe:

mihi sermo placuisset misi e uobis duodecim uiros singulos

1:24 And they turned and went up into the mountain, and de tribubus suis 1:24 qui cum perrexissent et ascendissent

came unto the valley of Eshcol, and searched it out.

in montana uenerunt usque ad uallem Botri et considerata

1:25 And they took of the fruit of the land in their hands, terra 1:25 sumentes de fructibus eius ut ostenderent uber-

and brought it down unto us, and brought us word again, and said, It is a good land which the LORD our God doth

give us. tatem adtulerunt ad nos atque dixerunt bona est terra quam 1:26 Notwithstanding ye would not go up, but rebelled Dominus Deus noster daturus est nobis 1:26 et noluistis as-

against the commandment of the LORD your God:

1:27 And ye murmured in your tents, and said, Because cendere sed increduli ad sermonem Domini Dei nostri 1:27

the LORD hated us, he hath brought us forth out of the land of Egypt, to deliver us into the hand of the Amorites,

to destroy us. murmurati estis in tabernaculis uestris atque dixistis odit nos

Dominus et idcirco eduxit nos de terra Aegypti ut traderet

1:28 Whither shall we go up? our brethren have discour- in manu Amorrei atque deleret 1:28 quo ascendemus nunaged our heart, saying, The people is greater and taller than we; the cities are great and walled up to heaven; and

moreover we have seen the sons of the Anakims there. tii terruerunt cor nostrum dicentes maxima multitudo est et

nobis in statura procerior urbes magnae et ad caelum usque

1:29 Then I said unto you, Dread not, neither be afraid of munitae filios Enacim uidimus ibi 1:29 et dixi uobis nothem.

1:30 The LORD your God which goeth before you, he lite metuere nec timeatis eos 1:30 Dominus Deus qui duc-

shall fight for you, according to all that he did for you in Egypt before your eyes;

tor est uester pro uobis ipse pugnabit sicut fecit in Aegypto

1:31 And in the wilderness, where thou hast seen how that uidentibus cunctis 1:31 et in solitudine ipse uidisti portauit

the LORD thy God bare thee, as a man doth bear his son, in all the way that ye went, until ye came into this place.

te Dominus Deus tuus ut solet homo gestare paruulum filium suum in omni uia per quam ambulasti donec ueniretis

1:32 Yet in this thing ye did not believe the LORD your ad locum istum 1:32 et nec sic quidem credidistis Domino

God,

Liber Deuteronomii

Deo uestro 1:33 qui praecessit uos in uia et metatus est lo- 1:33 Who went in the way before you, to search you out

a place to pitch your tents in, in fire by night, to shew you by what way ye should go, and in a cloud by day.

cum in quo tentoria figere deberetis nocte ostendens uobis iter per ignem et die per columnam nubis 1:34 cumque au- 1:34 And the LORD heard the voice of your words, and

was wroth, and sware, saying,

disset Dominus uocem sermonum uestrorum iratus iurauit et ait 1:35 non uidebit quispiam de hominibus generationis 1:35 Surely there shall not one of these men of this evil

generation see that good land, which I sware to give unto your fathers.

huius pessimae terram bonam quam sub iuramento pollicitus sum patribus uestris 1:36 praeter Chaleb filium Iepphonne 1:36 Save Caleb the son of Jephunneh; he shall see it, and

to him will I give the land that he hath trodden upon, and to his children, because he hath wholly followed the LORD.

ipse enim uidebit eam et ipsi dabo terram quam calcauit et filiis eius quia secutus est Dominum 1:37 nec miranda indi- 1:37 Also the LORD was angry with me for your sakes,

saying, Thou also shalt not go in thither.

gnatio in populum cum mihi quoque iratus Dominus propter uos dixerit nec tu ingredieris illuc 1:38 sed Iosue filius Nun 1:38 But Joshua the son of Nun, which standeth before

thee, he shall go in thither: encourage him: for he shall cause Israel to inherit it.

minister tuus ipse intrabit pro te hunc exhortare et robora et ipse terram sorte diuidat Israheli 1:39 paruuli uestri de qui- 1:39 Moreover your little ones, which ye said should be a

prey, and your children, which in that day had no knowledge between good and evil, they shall go in thither, and

bus dixistis quod captiui ducerentur et filii qui hodie boni unto them will I give it, and they shall possess it. ac mali ignorant distantiam ipsi ingredientur et ipsis dabo terram et possidebunt eam 1:40 uos autem reuertimini et 1:40 But as for you, turn you, and take your journey into

the wilderness by the way of the Red sea.

abite in solitudinem per uiam maris Rubri 1:41 et respon- 1:41 Then ye answered and said unto me, We have sinned

against the LORD, we will go up and fight, according to all that the LORD our God commanded us. And when ye had girded on every man his weapons of war, ye were ready to go up into the hill.

distis mihi peccauimus Domino ascendemus atque pugnabimus sicut praecepit Dominus Deus noster cumque instructi armis pergeretis in montem 1:42 ait mihi Dominus dic ad 1:42 And the LORD said unto me, Say unto them. Go

not up, neither fight; for I am not among you; lest ye be smitten before your enemies.

eos nolite ascendere neque pugnetis non enim sum uobiscum ne cadatis coram inimicis uestris 1:43 locutus sum et non 1:43 So I spake unto you; and ye would not hear, but re-

belled against the commandment of the LORD, and went presumptuously up into the hill.

audistis sed aduersantes imperio Domini et tumentes superbia ascendistis in montem 1:44 itaque egressus Amorreus 1:44 And the Amorites, which dwelt in that mountain,

came out against you, and chased you, as bees do, and destroyed you in Seir, even unto Hormah.

qui habitabat in montibus et obuiam ueniens persecutus est uos sicut solent apes persequi et cecidit de Seir usque Horma 1:45 cumque reuersi ploraretis coram Domino non audiuit 1:45 And ye returned and wept before the LORD; but the

LORD would not hearken to your voice, nor give ear unto you.

uos nec uoci uestrae uoluit adquiescere 1:46 sedistis ergo

1:46 So ye abode in Kadesh many days, according unto the days that ye abode there.

in Cadesbarne multo tempore

2:1 profectique inde uenimus in solitudinem quae ducit ad 2:1 Then we turned, and took our journey into the wilder-

ness by the way of the Red sea, as the LORD spake unto me: and we compassed mount Seir many days.

mare Rubrum sicut mihi dixerat Dominus et circumiuimus montem Seir longo tempore 2:2 dixitque Dominus ad me 2:2 And the LORD spake unto me, saying, 2:3 sufficit uobis circumire montem istum ite contra aquilo- 2:3 Ye have compassed this mountain long enough: turn

you northward.

Biblia Sacra Vulgata

2:4 And command thou the people, saying, Ye are to pass nem 2:4 et populo praecipe dicens transibitis per terminos

through the coast of your brethren the children of Esau, which dwell in Seir; and they shall be afraid of you: take

ye good heed unto yourselves therefore: fratrum uestrorum filiorum Esau qui habitant in Seir et ti-2:5 Meddle not with them; for I will not give you of their mebunt uos 2:5 uidete ergo diligenter ne moueamini contra

land, no, not so much as a foot breadth; because I have given mount Seir unto Esau for a possession.

eos neque enim dabo uobis de terra eorum quantum potest unius pedis calcare uestigium quia in possessionem Esau de-

2:6 Ye shall buy meat of them for money, that ye may eat; di montem Seir 2:6 cibos emetis ab eis pecunia et comeand ye shall also buy water of them for money, that ye may drink.

detis aquam emptam haurietis et bibetis 2:7 Dominus De-

2:7 For the LORD thy God hath blessed thee in all the works of thy hand: he knoweth thy walking through this great wilderness: these forty years the LORD thy God hath been with thee; thou hast lacked nothing.

us tuus benedixit tibi in omni opere manuum tuarum nouit iter tuum quomodo transieris solitudinem hanc magnam per quadraginta annos habitans tecum Dominus Deus tuus et ni-

2:8 And when we passed by from our brethren the children hil tibi defuit 2:8 cumque transissemus fratres nostros filios

of Esau, which dwelt in Seir, through the way of the plain from Elath, and from Eziongaber, we turned and passed

by the way of the wilderness of Moab. Esau qui habitabant in Seir per uiam campestrem de Helath

et de Asiongaber uenimus ad iter quod ducit in desertum

2:9 And the LORD said unto me, Distress not the Moab 2:9 dixitque Dominus ad me non pugnes contra Mo-

Moabites, neither contend with them in battle: for I will not give thee of their land for a possession; because I have

given Ar unto the children of Lot for a possession. abitas nec ineas aduersum eos proelium non enim dabo tibi

quicquam de terra eorum quia filiis Loth tradidi Ar in posses-

2:10 The Emims dwelt therein in times past, a people sionem 2:10 Emim primi fuerunt habitatores eius populus

great, and many, and tall, as the Anakims;

2:11 Which also were accounted giants, as the Anakims; magnus et ualidus et tam excelsus ut de Enacim stirpe 2:11

but the Moabites called them Emims.

quasi gigantes crederentur et essent similes filiorum Enacim

2:12 The Horims also dwelt in Seir beforetime; but the denique Moabitae appellant eos Emim 2:12 in Seir autem

children of Esau succeeded them, when they had destroyed them from before them, and dwelt in their stead; as Israel did unto the land of his possession, which the LORD gave unto them.

prius habitauerunt Horim quibus expulsis atque deletis habitauerunt filii Esau sicut fecit Israhel in terra possessionis

2:13 Now rise up, said I, and get you over the brook Zered. suae quam dedit ei Dominus 2:13 surgentes ergo ut trans-And we went over the brook Zered. 2:14 And the space in which we came from Kadeshbarnea, iremus torrentem Zared uenimus ad eum 2:14 tempus auuntil we were come over the brook Zered, was thirty and eight years; until all the generation of the men of war were wasted out from among the host, as the LORD sware unto them.

tem quo ambulauimus de Cadesbarne usque ad transitum torrentis Zared triginta octo annorum fuit donec consumeretur omnis generatio hominum bellatorum de castris sicut iu-

2:15 For indeed the hand of the LORD was against them, rauerat Dominus 2:15 cuius manus fuit aduersum eos ut in-

to destroy them from among the host, until they were consumed.

terirent de castrorum medio 2:16 postquam autem uniuersi

2:16 So it came to pass, when all the men of war were consumed and dead from among the people,

2:17 That the LORD spake unto me, saying, ceciderunt pugnatores 2:17 locutus est Dominus ad me di-2:18 Thou art to pass over through Ar, the coast of Moab, cens 2:18 tu transibis hodie terminos Moab urbem nomine

this day: 2:19 And when thou comest nigh over against the children Ar 2:19 et accedens in uicina filiorum Ammon caue ne pugof Ammon, distress them not, nor meddle with them: for I will not give thee of the land of the children of Ammon any possession; because I have given it unto the children of Lot for a possession.

nes contra eos nec mouearis ad proelium non enim dabo tibi

Liber Deuteronomii

de terra filiorum Ammon quia filiis Loth dedi eam in possessionem 2:20 terra gigantum reputata est et in ipsa olim 2:20 (That also was accounted a land of giants: giants

dwelt therein in old time; and the Ammonites call them Zamzummims;

habitauerunt gigantes quos Ammanitae uocant Zomzommim 2:21 populus magnus et multus et procerae longitudinis sic- 2:21 A people great, and many, and tall, as the Anakims;

but the LORD destroyed them before them; and they succeeded them, and dwelt in their stead:

ut Enacim quos deleuit Dominus a facie eorum et fecit illos habitare pro eis 2:22 sicut fecerat filiis Esau qui habitant 2:22 As he did to the children of Esau, which dwelt in Seir,

when he destroyed the Horims from before them; and they succeeded them, and dwelt in their stead even unto this

in Seir delens Horreos et terram eorum illis tradens quam day: possident usque in praesens 2:23 Eueos quoque qui habi- 2:23 And the Avims which dwelt in Hazerim, even unto

Azzah, the Caphtorims, which came forth out of Caphtor, destroyed them, and dwelt in their stead.)

tabant in Aserim usque Gazam Cappadoces expulerunt qui egressi de Cappadocia deleuerunt eos et habitauerunt pro illis 2:24 surgite et transite torrentem Arnon ecce tradidi in 2:24 Rise ye up, take your journey, and pass over the river

Arnon: behold, I have given into thine hand Sihon the Amorite, king of Heshbon, and his land: begin to possess

manu tua Seon regem Esebon Amorreum et terram eius in- it, and contend with him in battle. cipe possidere et committe aduersum eum proelium 2:25 2:25 This day will I begin to put the dread of thee and

the fear of thee upon the nations that are under the whole heaven, who shall hear report of thee, and shall tremble,

hodie incipiam mittere terrorem atque formidinem tuam in and be in anguish because of thee. populos qui habitant sub omni caelo ut audito nomine tuo paueant et in morem parturientium contremescant et dolore teneantur 2:26 misi ergo nuntios de solitudine Cademoth 2:26 And I sent messengers out of the wilderness of Kede-

moth unto Sihon king of Heshbon with words of peace, saying,

ad Seon regem Esebon uerbis pacificis dicens 2:27 trans-

2:27 Let me pass through thy land: I will go along by the high way, I will neither turn unto the right hand nor to the left.

ibimus per terram tuam publica gradiemur uia non declinabimus neque ad dextram neque ad sinistram 2:28 alimenta 2:28 Thou shalt sell me meat for money, that I may eat;

and give me water for money, that I may drink: only I will pass through on my feet;

pretio uende nobis ut uescamur aquam pecunia tribue et sic bibemus tantum est ut nobis concedas transitum 2:29 sicut 2:29 (As the children of Esau which dwell in Seir, and the

Moabites which dwell in Ar, did unto me;) until I shall pass over Jordan into the land which the LORD our God

fecerunt filii Esau qui habitant in Seir et Moabitae qui mo- giveth us. rantur in Ar donec ueniamus ad Iordanem et transeamus in terram quam Dominus Deus noster daturus est nobis 2:30 2:30 But Sihon king of Heshbon would not let us pass by

him: for the LORD thy God hardened his spirit, and made his heart obstinate, that he might deliver him into thy hand,

noluitque Seon rex Esebon dare nobis transitum quia indu- as appeareth this day. rauerat Dominus Deus tuus spiritum eius et obfirmauerat cor illius ut traderetur in manus tuas sicut nunc uides 2:31 dixit- 2:31 And the LORD said unto me, Behold, I have begun

to give Sihon and his land before thee: begin to possess, that thou mayest inherit his land.

que Dominus ad me ecce coepi tradere tibi Seon et terram

eius incipe possidere eam 2:32 egressusque est Seon ob- 2:32 Then Sihon came out against us, he and all his people,

to fight at Jahaz.

uiam nobis cum omni populo suo ad proelium in Iesa 2:33 2:33 And the LORD our God delivered him before us; and

we smote him, and his sons, and all his people.

et tradidit eum Dominus Deus noster nobis percussimusque

eum cum filiis et omni populo suo 2:34 cunctasque urbes 2:34 And we took all his cities at that time, and utterly destroyed the men, and the women, and the little ones, of every city, we left none to remain:

Biblia Sacra Vulgata

in tempore illo cepimus interfectis habitatoribus earum uiris ac mulieribus et paruulis non reliquimus in eis quicquam

2:35 Only the cattle we took for a prey unto ourselves, and 2:35 absque iumentis quae in partem uenere praedantium et

the spoil of the cities which we took.

2:36 From Aroer, which is by the brink of the river of spoliis urbium quas cepimus 2:36 ab Aroer quae est super

Arnon, and from the city that is by the river, even unto Gilead, there was not one city too strong for us: the LORD

our God delivered all unto us: ripam torrentis Arnon oppido quod in ualle situm est usque

Galaad non fuit uicus et ciuitas quae nostras effugeret ma-

2:37 Only unto the land of the children of Ammon thou nus omnes tradidit Dominus Deus noster nobis 2:37 absque

camest not, nor unto any place of the river Jabbok, nor unto the cities in the mountains, nor unto whatsoever the

LORD our God forbad us. terra filiorum Ammon ad quam non accessimus et cunctis quae adiacent torrenti Ieboc et urbibus montanis uniuersisque locis a quibus nos prohibuit Dominus Deus noster

3:1 Then we turned, and went up the way to Bashan: and 3:1 itaque conuersi ascendimus per iter Basan egressus-

Og the king of Bashan came out against us, he and all his people, to battle at Edrei.

que est Og rex Basan in occursum nobis cum populo suo ad

3:2 And the LORD said unto me, Fear him not: for I will bellandum in Edrai 3:2 dixitque Dominus ad me ne timeas

deliver him, and all his people, and his land, into thy hand; and thou shalt do unto him as thou didst unto Sihon king

of the Amorites, which dwelt at Heshbon. eum quia in manu tua traditus est cum omni populo ac terra

sua faciesque ei sicut fecisti Seon regi Amorreorum qui ha-

3:3 So the LORD our God delivered into our hands Og bitauit in Esebon 3:3 tradidit ergo Dominus Deus noster in

also, the king of Bashan, and all his people: and we smote him until none was left to him remaining.

manibus nostris etiam Og regem Basan et uniuersum popu-

3:4 And we took all his cities at that time, there was not lum eius percussimusque eos usque ad internicionem 3:4

a city which we took not from them, threescore cities, all the region of Argob, the kingdom of Og in Bashan.

uastantes cunctas ciuitates illius uno tempore non fuit oppidum quod nos effugeret sexaginta urbes omnem regionem

3:5 All these cities were fenced with high walls, gates, and Argob regni Og in Basan 3:5 cunctae urbes erant munitae bars; beside unwalled towns a great many.

muris altissimis portisque et uectibus absque oppidis innu-

3:6 And we utterly destroyed them, as we did unto Sihon meris quae non habebant muros 3:6 et deleuimus eos sicut

king of Heshbon, utterly destroying the men, women, and children, of every city.

feceramus Seon regi Esebon disperdentes omnem ciuitatem

3:7 But all the cattle, and the spoil of the cities, we took uirosque ac mulieres et paruulos 3:7 iumenta autem et spo-

for a prey to ourselves.

3:8 And we took at that time out of the hand of the two lia urbium diripuimus 3:8 tulimusque illo in tempore ter-

kings of the Amorites the land that was on this side Jordan, from the river of Arnon unto mount Hermon;

ram de manu duorum regum Amorreorum qui erant trans

3:9 (Which Hermon the Sidonians call Sirion; and the Iordanem a torrente Arnon usque ad montem Hermon 3:9 Amorites call it Shenir;) 3:10 All the cities of the plain, and all Gilead, and all quem Sidonii Sarion uocant et Amorrei Sanir 3:10 omnes Bashan, unto Salchah and Edrei, cities of the kingdom of Og in Bashan.

ciuitates quae sitae sunt in planitie et uniuersam terram Galaad et Basan usque Selcha et Edrai ciuitates regni Og in

3:11 For only Og king of Bashan remained of the remnant Basan 3:11 solus quippe Og rex Basan restiterat de stirpe

of giants; behold his bedstead was a bedstead of iron; is it not in Rabbath of the children of Ammon? nine cubits was the length thereof, and four cubits the breadth of it, after the cubit of a man.

gigantum monstratur lectus eius ferreus qui est in Rabbath

Liber Deuteronomii

filiorum Ammon nouem cubitos habens longitudinis et quattuor latitudinis ad mensuram cubiti uirilis manus 3:12 ter- 3:12 And this land, which we possessed at that time, from

Aroer, which is by the river Arnon, and half mount Gilead, and the cities thereof, gave I unto the Reubenites and to the

ramque possedimus in tempore illo ab Aroer quae est super Gadites. ripam torrentis Arnon usque ad mediam partem montis Galaad et ciuitates illius dedi Ruben et Gad 3:13 reliquam au- 3:13 And the rest of Gilead, and all Bashan, being the

kingdom of Og, gave I unto the half tribe of Manasseh; all the region of Argob, with all Bashan, which was called

tem partem Galaad et omnem Basan regni Og tradidi mediae the land of giants. tribui Manasse omnem regionem Argob cuncta Basan uocatur terra gigantum 3:14 Iair filius Manasse possedit omnem 3:14 Jair the son of Manasseh took all the country of Ar-

gob unto the coasts of Geshuri and Maachathi; and called them after his own name, Bashanhavothjair, unto this day.

regionem Argob usque ad terminos Gesuri et Machathi uocauitque ex nomine suo Basan Auothiair id est uillas Iair usque in praesentem diem 3:15 Machir quoque dedi Ga- 3:15 And I gave Gilead unto Machir. laad 3:16 et tribubus Ruben et Gad dedi terram Galaad us- 3:16 And unto the Reubenites and unto the Gadites I gave

from Gilead even unto the river Arnon half the valley, and the border even unto the river Jabbok, which is the border

que ad torrentem Arnon medium torrentis et finium usque of the children of Ammon; ad torrentem Ieboc qui est terminus filiorum Ammon 3:17 3:17 The plain also, and Jordan, and the coast thereof,

from Chinnereth even unto the sea of the plain, even the salt sea, under Ashdothpisgah eastward.

et planitiem solitudinis atque Iordanem et terminos Chenereth usque ad mare Deserti quod est Salsissimum ad radices montis Phasga contra orientem 3:18 praecepique uobis in 3:18 And I commanded you at that time, saying, The

LORD your God hath given you this land to possess it: ye shall pass over armed before your brethren the children

tempore illo dicens Dominus Deus uester dat uobis terram of Israel, all that are meet for the war. hanc in hereditatem expediti praecedite fratres uestros filios Israhel omnes uiri robusti 3:19 absque uxoribus et paruulis 3:19 But your wives, and your little ones, and your cattle,

(for I know that ye have much cattle,) shall abide in your cities which I have given you;

ac iumentis noui enim quod plura habeatis pecora et in urbibus remanere debebunt quas tradidi uobis 3:20 donec re- 3:20 Until the LORD have given rest unto your brethren,

as well as unto you, and until they also possess the land which the LORD your God hath given them beyond Jordan: and then shall ye return every man unto his possession, which I have given you.

quiem tribuat Dominus fratribus uestris sicut uobis tribuit et possideant etiam ipsi terram quam daturus est eis trans Iordanem tunc reuertetur unusquisque in possessionem suam quam dedi uobis 3:21 Iosue quoque in tempore illo prae- 3:21 And I commanded Joshua at that time, saying, Thine

eyes have seen all that the LORD your God hath done unto these two kings: so shall the LORD do unto all the king-

cepi dicens oculi tui uiderunt quae fecit Dominus Deus ue- doms whither thou passest. ster duobus his regibus sic faciet omnibus regnis ad quae transiturus es 3:22 ne timeas eos Dominus enim Deus ues- 3:22 Ye shall not fear them: for the LORD your God he

shall fight for you.

ter pugnabit pro uobis 3:23 precatusque sum Dominum in 3:23 And I besought the LORD at that time, saying, tempore illo dicens 3:24 Domine Deus tu coepisti osten- 3:24 O Lord GOD, thou hast begun to shew thy servant

thy greatness, and thy mighty hand: for what God is there in heaven or in earth, that can do according to thy works,

dere seruo tuo magnitudinem tuam manumque fortissimam and according to thy might? neque enim est alius Deus uel in caelo uel in terra qui pos-

Biblia Sacra Vulgata

3:25 I pray thee, let me go over, and see the good land that sit facere opera tua et conparari fortitudini tuae 3:25 trans-

is beyond Jordan, that goodly mountain, and Lebanon.

ibo igitur et uidebo terram hanc optimam trans Iordanem et

3:26 But the LORD was wroth with me for your sakes, montem istum egregium et Libanum 3:26 iratusque est Do-

and would not hear me: and the LORD said unto me, Let it suffice thee; speak no more unto me of this matter.

minus mihi propter uos nec exaudiuit me sed dixit mihi suf-

3:27 Get thee up into the top of Pisgah, and lift up thine ficit tibi nequaquam ultra loquaris de hac re ad me 3:27 as-

eyes westward, and northward, and southward, and eastward, and behold it with thine eyes: for thou shalt not go

over this Jordan. cende cacumen Phasgae et oculos tuos circumfer ad occiden-

tem et aquilonem austrumque et orientem et aspice nec enim

3:28 But charge Joshua, and encourage him, and transibis Iordanem istum 3:28 praecipe Iosue et corrobora

strengthen him: for he shall go over before this people, and he shall cause them to inherit the land which thou shalt

see. eum atque conforta quia ipse praecedet populum istum et di-3:29 So we abode in the valley over against Bethpeor. uidet eis terram quam uisurus es 3:29 mansimusque in ualle

contra fanum Phogor

4:1 Now therefore hearken, O Israel, unto the statutes and 4:1 et nunc Israhel audi praecepta et iudicia quae ego dounto the judgments, which I teach you, for to do them, that ye may live, and go in and possess the land which the

LORD God of your fathers giveth you. ceo te ut faciens ea uiuas et ingrediens possideas terram

quam Dominus Deus patrum uestrorum daturus est uobis

4:2 Ye shall not add unto the word which I command you, 4:2 non addetis ad uerbum quod uobis loquor neque auferetis

neither shall ye diminish ought from it, that ye may keep the commandments of the LORD your God which I com-

mand you. ex eo custodite mandata Domini Dei uestri quae ego praeci-4:3 Your eyes have seen what the LORD did because pio uobis 4:3 oculi uestri uiderunt omnia quae fecit Domi-

of Baalpeor: for all the men that followed Baalpeor, the LORD thy God hath destroyed them from among you.

nus contra Beelphegor quomodo contriuerit omnes cultores

4:4 But ye that did cleave unto the LORD your God are eius de medio uestri 4:4 uos autem qui adheretis Domino

alive every one of you this day. 4:5 Behold, I have taught you statutes and judgments, even Deo uestro uiuitis uniuersi usque in praesentem diem 4:5 as the LORD my God commanded me, that ye should do so in the land whither ye go to possess it.

scitis quod docuerim uos praecepta atque iustitias sicut mandauit mihi Dominus Deus meus sic facietis ea in terra quam

4:6 Keep therefore and do them; for this is your wisdom possessuri estis 4:6 et obseruabitis et implebitis opere haec

and your understanding in the sight of the nations, which shall hear all these statutes, and say, Surely this great na-

tion is a wise and understanding people. est enim uestra sapientia et intellectus coram populis ut au-

dientes uniuersa praecepta haec dicant en populus sapiens et

4:7 For what nation is there so great, who hath God so intellegens gens magna 4:7 nec est alia natio tam grandis

nigh unto them, as the LORD our God is in all things that we call upon him for?

quae habeat deos adpropinquantes sibi sicut Dominus De-

4:8 And what nation is there so great, that hath statutes and us noster adest cunctis obsecrationibus nostris 4:8 quae est

judgments so righteous as all this law, which I set before you this day?

enim alia gens sic inclita ut habeat caerimonias iustaque iudicia et uniuersam legem quam ego proponam hodie ante

4:9 Only take heed to thyself, and keep thy soul diligently, oculos uestros 4:9 custodi igitur temet ipsum et animam lest thou forget the things which thine eyes have seen, and lest they depart from thy heart all the days of thy life: but

teach them thy sons, and thy sons’ sons; tuam sollicite ne obliuiscaris uerborum quae uiderunt oculi tui et ne excedant de corde tuo cunctis diebus uitae tuae

Liber Deuteronomii

docebis ea filios ac nepotes tuos 4:10 diem in quo stetisti 4:10 Specially the day that thou stoodest before the LORD

thy God in Horeb, when the LORD said unto me, Gather me the people together, and I will make them hear my words, that they may learn to fear me all the days that they shall live upon the earth, and that they may teach their children.

coram Domino Deo tuo in Horeb quando Dominus locutus est mihi dicens congrega ad me populum ut audiat sermones meos et discat timere me omni tempore quo uiuit in terra doceantque filios suos 4:11 et accessistis ad radices montis 4:11 And ye came near and stood under the mountain; and

the mountain burned with fire unto the midst of heaven, with darkness, clouds, and thick darkness.

qui ardebat usque ad caelum erantque in eo tenebrae nubes et caligo 4:12 locutusque est Dominus ad uos de medio ignis 4:12 And the LORD spake unto you out of the midst of

the fire: ye heard the voice of the words, but saw no similitude; only ye heard a voice.

uocem uerborum eius audistis et formam penitus non uidistis 4:13 et ostendit uobis pactum suum quod praecepit ut 4:13 And he declared unto you his covenant, which he

commanded you to perform, even ten commandments; and he wrote them upon two tables of stone.

faceretis et decem uerba quae scripsit in duabus tabulis lapideis 4:14 mihique mandauit in illo tempore ut docerem uos 4:14 And the LORD commanded me at that time to teach

you statutes and judgments, that ye might do them in the land whither ye go over to possess it.

caerimonias et iudicia quae facere deberetis in terra quam possessuri estis 4:15 custodite igitur sollicite animas ue- 4:15 Take ye therefore good heed unto yourselves; for ye

saw no manner of similitude on the day that the LORD spake unto you in Horeb out of the midst of the fire:

stras non uidistis aliquam similitudinem in die qua locutus est Dominus uobis in Horeb de medio ignis 4:16 ne forte 4:16 Lest ye corrupt yourselves, and make you a graven

image, the similitude of any figure, the likeness of male or female,

decepti faciatis uobis sculptam similitudinem aut imaginem masculi uel feminae 4:17 similitudinem omnium iumento- 4:17 The likeness of any beast that is on the earth, the

likeness of any winged fowl that flieth in the air,

rum quae sunt super terram uel auium sub caelo uolantium 4:18 atque reptilium quae mouentur in terra siue piscium qui 4:18 The likeness of any thing that creepeth on the ground,

the likeness of any fish that is in the waters beneath the earth:

sub terra morantur in aquis 4:19 ne forte oculis eleuatis ad

4:19 And lest thou lift up thine eyes unto heaven, and when thou seest the sun, and the moon, and the stars, even all the host of heaven, shouldest be driven to worship them, and serve them, which the LORD thy God hath divided unto all nations under the whole heaven.

caelum uideas solem et lunam et omnia astra caeli et errore deceptus adores ea et colas quae creauit Dominus Deus tuus in ministerium cunctis gentibus quae sub caelo sunt 4:20 4:20 But the LORD hath taken you, and brought you forth

out of the iron furnace, even out of Egypt, to be unto him a people of inheritance, as ye are this day.

uos autem tulit Dominus et eduxit de fornace ferrea Aegypti ut haberet populum hereditarium sicut est in praesenti die 4:21 iratusque est Dominus contra me propter sermones 4:21 Furthermore the LORD was angry with me for your

sakes, and sware that I should not go over Jordan, and that I should not go in unto that good land, which the LORD

uestros et iurauit ut non transirem Iordanem nec ingrede- thy God giveth thee for an inheritance:

rer terram optimam quam daturus est uobis 4:22 ecce mo- 4:22 But I must die in this land, I must not go over Jordan:

but ye shall go over, and possess that good land.

rior in hac humo non transibo Iordanem uos transibitis et possidebitis terram egregiam 4:23 caue nequando obliuis- 4:23 Take heed unto yourselves, lest ye forget the covenant

of the LORD your God, which he made with you, and make you a graven image, or the likeness of any thing,

caris pacti Domini Dei tui quod pepigit tecum et facias tibi which the LORD thy God hath forbidden thee. sculptam similitudinem eorum quae fieri Dominus prohibuit 4:24 quia Dominus Deus tuus ignis consumens est Deus ae- 4:24 For the LORD thy God is a consuming fire, even a

jealous God.

Biblia Sacra Vulgata

4:25 When thou shalt beget children, and children’s chil- mulator 4:25 si genueritis filios ac nepotes et morati fueritis

dren, and ye shall have remained long in the land, and shall corrupt yourselves, and make a graven image, or the likeness of any thing, and shall do evil in the sight of the LORD thy God, to provoke him to anger:

in terra deceptique feceritis uobis aliquam similitudinem patrantes malum coram Domino Deo uestro ut eum ad iracun-

4:26 I call heaven and earth to witness against you this day, diam prouocetis 4:26 testes inuoco hodie caelum et terram that ye shall soon utterly perish from off the land whereunto ye go over Jordan to possess it; ye shall not prolong

your days upon it, but shall utterly be destroyed. cito perituros uos esse de terra quam transito Iordane pos-

sessuri estis non habitabitis in ea longo tempore sed delebit

4:27 And the LORD shall scatter you among the nations, uos Dominus 4:27 atque disperget in omnes gentes et rema-

and ye shall be left few in number among the heathen, whither the LORD shall lead you.

nebitis pauci in nationibus ad quas uos ducturus est Domi-

4:28 And there ye shall serve gods, the work of men’s nus 4:28 ibique seruietis diis qui hominum manu fabricati

hands, wood and stone, which neither see, nor hear, nor eat, nor smell.

sunt ligno et lapidi qui non uident non audiunt non comedunt

4:29 But if from thence thou shalt seek the LORD thy God, non odorantur 4:29 cumque quaesieris ibi Dominum Deum thou shalt find him, if thou seek him with all thy heart and with all thy soul.

tuum inuenies eum si tamen toto corde quaesieris et tota tri-

4:30 When thou art in tribulation, and all these things are bulatione animae tuae 4:30 postquam te inuenerint omnia come upon thee, even in the latter days, if thou turn to the LORD thy God, and shalt be obedient unto his voice;

quae praedicta sunt nouissimo tempore reuerteris ad Domi-

4:31 (For the LORD thy God is a merciful God;) he num Deum tuum et audies uocem eius 4:31 quia Deus mi-

will not forsake thee, neither destroy thee, nor forget the covenant of thy fathers which he sware unto them.

sericors Dominus Deus tuus est non dimittet te nec omnino delebit neque obliuiscetur pacti in quo iurauit patribus tuis

4:32 For ask now of the days that are past, which were 4:32 interroga de diebus antiquis qui fuerunt ante te ex die

before thee, since the day that God created man upon the earth, and ask from the one side of heaven unto the other, whether there hath been any such thing as this great thing is, or hath been heard like it?

quo creauit Deus hominem super terram a summo caeli usque ad summum eius si facta est aliquando huiuscemodi res

4:33 Did ever people hear the voice of God speaking out aut umquam cognitum est 4:33 ut audiret populus uocem of the midst of the fire, as thou hast heard, and live?

4:34 Or hath God assayed to go and take him a nation from Dei loquentis de medio ignis sicut tu audisti et uixisti 4:34

the midst of another nation, by temptations, by signs, and by wonders, and by war, and by a mighty hand, and by a stretched out arm, and by great terrors, according to all that the LORD your God did for you in Egypt before your eyes?

si fecit Deus ut ingrederetur et tolleret sibi gentem de medio nationum per temptationes signa atque portenta per pugnam et robustam manum extentumque brachium et horribiles uisiones iuxta omnia quae fecit pro uobis Dominus Deus ues-

4:35 Unto thee it was shewed, that thou mightest know ter in Aegypto uidentibus oculis tuis 4:35 ut scires quoniam

that the LORD he is God; there is none else beside him.

4:36 Out of heaven he made thee to hear his voice, that Dominus ipse est Deus et non est alius praeter unum 4:36

he might instruct thee: and upon earth he shewed thee his great fire; and thou heardest his words out of the midst of

the fire. de caelo te fecit audire uocem suam ut doceret te et in terra

ostendit tibi ignem suum maximum et audisti uerba illius de

4:37 And because he loved thy fathers, therefore he chose medio ignis 4:37 quia dilexit patres tuos et elegit semen eo-

their seed after them, and brought thee out in his sight with his mighty power out of Egypt;

rum post eos eduxitque te praecedens in uirtute sua magna

4:38 To drive out nations from before thee greater and ex Aegypto 4:38 ut deleret nationes maximas et fortiores te

mightier than thou art, to bring thee in, to give thee their land for an inheritance, as it is this day.

Liber Deuteronomii

in introitu tuo et introduceret te daretque tibi terram earum in possessionem sicut cernis in praesenti die 4:39 scito ergo 4:39 Know therefore this day, and consider it in thine

heart, that the LORD he is God in heaven above, and upon the earth beneath: there is none else.

hodie et cogitato in corde tuo quod Dominus ipse sit Deus in caelo sursum et in terra deorsum et non sit alius 4:40 cus- 4:40 Thou shalt keep therefore his statutes, and his com-

mandments, which I command thee this day, that it may go well with thee, and with thy children after thee, and that thou mayest prolong thy days upon the earth, which the LORD thy God giveth thee, for ever.

todi praecepta eius atque mandata quae ego praecipio tibi ut bene sit tibi et filiis tuis post te et permaneas multo tempore super terram quam Dominus Deus tuus daturus est tibi 4:41 4:41 Then Moses severed three cities on this side Jordan

toward the sunrising;

tunc separauit Moses tres ciuitates trans Iordanem ad orientalem plagam 4:42 ut confugiat ad eas qui occiderit nolens 4:42 That the slayer might flee thither, which should kill

his neighbour unawares, and hated him not in times past; and that fleeing unto one of these cities he might live:

proximum suum nec fuerit inimicus ante unum et alterum

diem et ad harum aliquam urbium possit euadere 4:43 Bo- 4:43 Namely, Bezer in the wilderness, in the plain country, of the Reubenites; and Ramoth in Gilead, of the Gadites; and Golan in Bashan, of the Manassites.

sor in solitudine quae sita est in terra campestri de tribu Ruben et Ramoth in Galaad quae est in tribu Gad et Golam in Basan quae est in tribu Manasse 4:44 ista est lex quam pro- 4:44 And this is the law which Moses set before the chil-

dren of Israel:

posuit Moses coram filiis Israhel 4:45 et haec testimonia 4:45 These are the testimonies, and the statutes, and the

judgments, which Moses spake unto the children of Israel, after they came forth out of Egypt.

et caerimoniae atque iudicia quae locutus est ad filios Israhel quando egressi sunt de Aegypto 4:46 trans Iordanem in 4:46 On this side Jordan, in the valley over against Beth-

peor, in the land of Sihon king of the Amorites, who dwelt at Heshbon, whom Moses and the children of Israel smote,

ualle contra fanum Phogor in terra Seon regis Amorrei qui after they were come forth out of Egypt: habitauit in Esebon quem percussit Moses filii quoque Israhel egressi ex Aegypto 4:47 possederunt terram eius et ter- 4:47 And they possessed his land, and the land of Og king

of Bashan, two kings of the Amorites, which were on this side Jordan toward the sunrising;

ram Og regis Basan duorum regum Amorreorum qui erant trans Iordanem ad solis ortum 4:48 ab Aroer quae sita est 4:48 From Aroer, which is by the bank of the river Arnon,

even unto mount Sion, which is Hermon,

super ripam torrentis Arnon usque ad montem Sion qui est et Hermon 4:49 omnem planitiem trans Iordanem ad orien- 4:49 And all the plain on this side Jordan eastward, even

unto the sea of the plain, under the springs of Pisgah.

talem plagam usque ad mare Solitudinis et usque ad radices montis Phasga

5:1 uocauitque Moses omnem Israhelem et dixit ad eum 5:1 And Moses called all Israel, and said unto them, Hear, O Israel, the statutes and judgments which I speak in your ears this day, that ye may learn them, and keep, and do

audi Israhel caerimonias atque iudicia quae ego loquor in them. auribus uestris hodie discite ea et opere conplete 5:2 Do- 5:2 The LORD our God made a covenant with us in Horeb. minus Deus noster pepigit nobiscum foedus in Horeb 5:3 5:3 The LORD made not this covenant with our fathers,

but with us, even us, who are all of us here alive this day.

non cum patribus nostris iniit pactum sed nobiscum qui inpraesentiarum sumus et uiuimus 5:4 facie ad faciem locu- 5:4 The LORD talked with you face to face in the mount

out of the midst of the fire,

tus est nobis in monte de medio ignis 5:5 ego sequester et 5:5 (I stood between the LORD and you at that time, to

shew you the word of the LORD: for ye were afraid by reason of the fire, and went not up into the mount;) saying,

Biblia Sacra Vulgata

medius fui inter Dominum et uos in tempore illo ut adnuntiarem uobis uerba eius timuistis enim ignem et non ascendistis

5:6 I am the LORD thy God, which brought thee out of in montem et ait 5:6 ego Dominus Deus tuus qui eduxi te

the land of Egypt, from the house of bondage.

5:7 Thou shalt have none other gods before me. de terra Aegypti de domo seruitutis 5:7 non habebis deos 5:8 Thou shalt not make thee any graven image, or any alienos in conspectu meo 5:8 non facies tibi sculptile nec

likeness of any thing that is in heaven above, or that is in the earth beneath, or that is in the waters beneath the earth:

similitudinem omnium quae in caelo sunt desuper et quae in

5:9 Thou shalt not bow down thyself unto them, nor serve terra deorsum et quae uersantur in aquis sub terra 5:9 non

them: for I the LORD thy God am a jealous God, visiting the iniquity of the fathers upon the children unto the third

and fourth generation of them that hate me, adorabis ea et non coles ego enim sum Dominus Deus tuus

Deus aemulator reddens iniquitatem patrum super filios in

5:10 And shewing mercy unto thousands of them that love tertiam et quartam generationem his qui oderunt me 5:10 et

me and keep my commandments.

faciens misericordiam in multa milia diligentibus me et cu-

5:11 Thou shalt not take the name of the LORD thy God in stodientibus praecepta mea 5:11 non usurpabis nomen Do-

vain: for the LORD will not hold him guiltless that taketh his name in vain.

mini Dei tui frustra quia non erit inpunitus qui super re uana

5:12 Keep the sabbath day to sanctify it, as the LORD thy nomen eius adsumpserit 5:12 obserua diem sabbati ut sanc-

God hath commanded thee.

5:13 Six days thou shalt labour, and do all thy work: tifices eum sicut praecepit tibi Dominus Deus tuus 5:13 sex 5:14 But the seventh day is the sabbath of the LORD thy diebus operaberis et facies omnia opera tua 5:14 septimus

God: in it thou shalt not do any work, thou, nor thy son, nor thy daughter, nor thy manservant, nor thy maidservant, nor thine ox, nor thine ass, nor any of thy cattle, nor thy stranger that is within thy gates; that thy manservant and thy maidservant may rest as well as thou.

dies sabbati est id est requies Domini Dei tui non facies in eo quicquam operis tu et filius tuus et filia seruus et ancilla et bos et asinus et omne iumentum tuum et peregrinus qui est intra portas tuas ut requiescat seruus et ancilla tua sic-

5:15 And remember that thou wast a servant in the land ut et tu 5:15 memento quod et ipse seruieris in Aegypto

of Egypt, and that the LORD thy God brought thee out thence through a mighty hand and by a stretched out arm: therefore the LORD thy God commanded thee to keep the sabbath day.

et eduxerit te inde Dominus Deus tuus in manu forti et brachio extento idcirco praecepit tibi ut obseruares diem sabbati

5:16 Honour thy father and thy mother, as the LORD thy 5:16 honora patrem tuum et matrem sicut praecepit tibi Do-

God hath commanded thee; that thy days may be prolonged, and that it may go well with thee, in the land which

the LORD thy God giveth thee. minus Deus tuus ut longo uiuas tempore et bene sit tibi in 5:17 Thou shalt not kill. terra quam Dominus Deus tuus daturus est tibi 5:17 non 5:18 Neither shalt thou commit adultery. occides 5:18 neque moechaberis 5:19 furtumque non fa-

5:19 Neither shalt thou steal.

5:20 Neither shalt thou bear false witness against thy cies 5:20 nec loqueris contra proximum tuum falsum testi-

neighbour. 5:21 Neither shalt thou desire thy neighbour’s wife, nei- monium 5:21 non concupisces uxorem proximi tui non dother shalt thou covet thy neighbour’s house, his field, or his manservant, or his maidservant, his ox, or his ass, or

any thing that is thy neighbour’s. mum non agrum non seruum non ancillam non bouem non 5:22 These words the LORD spake unto all your assembly asinum et uniuersa quae illius sunt 5:22 haec uerba locutus

in the mount out of the midst of the fire, of the cloud, and of the thick darkness, with a great voice: and he added no more. And he wrote them in two tables of stone, and delivered them unto me.

est Dominus ad omnem multitudinem uestram in monte de medio ignis et nubis et caliginis uoce magna nihil addens

Liber Deuteronomii

amplius et scripsit ea in duabus tabulis lapideis quas tradidit mihi 5:23 uos autem postquam audistis uocem de medio te- 5:23 And it came to pass, when ye heard the voice out of

the midst of the darkness, (for the mountain did burn with fire,) that ye came near unto me, even all the heads of your

nebrarum et montem ardere uidistis accessistis ad me omnes tribes, and your elders; principes tribuum et maiores natu atque dixistis 5:24 ecce 5:24 And ye said, Behold, the LORD our God hath shewed

us his glory and his greatness, and we have heard his voice out of the midst of the fire: we have seen this day that God

ostendit nobis Dominus Deus noster maiestatem et magnitu- doth talk with man, and he liveth. dinem suam uocem eius audiuimus de medio ignis et probauimus hodie quod loquente Deo cum homine uixerit homo 5:25 cur ergo morimur et deuorabit nos ignis hic maximus si 5:25 Now therefore why should we die? for this great fire

will consume us: if we hear the voice of the LORD our God any more, then we shall die.

enim audierimus ultra uocem Domini Dei nostri moriemur

5:26 quid est omnis caro ut audiat uocem Dei uiuentis qui 5:26 For who is there of all flesh, that hath heard the voice of the living God speaking out of the midst of the fire, as we have, and lived?

de medio ignis loquitur sicut nos audiuimus et possit uiuere 5:27 tu magis accede et audi cuncta quae dixerit Dominus 5:27 Go thou near, and hear all that the LORD our God

shall say: and speak thou unto us all that the LORD our God shall speak unto thee; and we will hear it, and do it.

Deus noster tibi loquerisque ad nos et nos audientes faciemus ea 5:28 quod cum audisset Dominus ait ad me audiui 5:28 And the LORD heard the voice of your words, when

ye spake unto me; and the LORD said unto me, I have heard the voice of the words of this people, which they have spoken unto thee: they have well said all that they have spoken.

uocem uerborum populi huius quae locuti sunt tibi bene omnia sunt locuti 5:29 quis det talem eos habere mentem ut 5:29 O that there were such an heart in them, that they

would fear me, and keep all my commandments always, that it might be well with them, and with their children for

timeant me et custodiant uniuersa mandata mea in omni tem- ever! pore ut bene sit eis et filiis eorum in sempiternum 5:30 uade 5:30 Go say to them, Get you into your tents again. et dic eis reuertimini in tentoria uestra 5:31 tu uero hic sta 5:31 But as for thee, stand thou here by me, and I will

speak unto thee all the commandments, and the statutes, and the judgments, which thou shalt teach them, that they

mecum et loquar tibi omnia mandata et caerimonias atque may do them in the land which I give them to possess it. iudicia quae docebis eos ut faciant ea in terra quam dabo illis in possessionem 5:32 custodite igitur et facite quae praece- 5:32 Ye shall observe to do therefore as the LORD your

God hath commanded you: ye shall not turn aside to the right hand or to the left.

pit Dominus Deus uobis non declinabitis neque ad dextram neque ad sinistram 5:33 sed per uiam quam praecepit Do- 5:33 Ye shall walk in all the ways which the LORD your

God hath commanded you, that ye may live, and that it may be well with you, and that ye may prolong your days

minus Deus uester ambulabitis ut uiuatis et bene sit uobis et in the land which ye shall possess. protelentur dies in terra possessionis uestrae

6:1 haec sunt praecepta et caerimoniae atque iudicia quae 6:1 Now these are the commandments, the statutes, and

the judgments, which the LORD your God commanded to teach you, that ye might do them in the land whither ye go

mandauit Dominus Deus uester ut docerem uos et faciatis ea to possess it: in terra ad quam transgredimini possidendam 6:2 ut timeas 6:2 That thou mightest fear the LORD thy God, to keep

all his statutes and his commandments, which I command thee, thou, and thy son, and thy son’s son, all the days of

Dominum Deum tuum et custodias omnia mandata et prae- thy life; and that thy days may be prolonged. cepta eius quae ego praecipio tibi et filiis ac nepotibus tuis cunctis diebus uitae tuae ut prolongentur dies tui 6:3 audi 6:3 Hear therefore, O Israel, and observe to do it; that it

may be well with thee, and that ye may increase mightily, as the LORD God of thy fathers hath promised thee, in the

Israhel et obserua ut facias et bene sit tibi et multipliceris land that floweth with milk and honey.

Biblia Sacra Vulgata

amplius sicut pollicitus est Dominus Deus patrum tuorum

6:4 Hear, O Israel: The LORD our God is one LORD: tibi terram lacte et melle manantem 6:4 audi Israhel Do-6:5 And thou shalt love the LORD thy God with all thine minus Deus noster Dominus unus est 6:5 diliges Dominum

heart, and with all thy soul, and with all thy might.

Deum tuum ex toto corde tuo et ex tota anima tua et ex tota

6:6 And these words, which I command thee this day, shall fortitudine tua 6:6 eruntque uerba haec quae ego praecipio

be in thine heart:

6:7 And thou shalt teach them diligently unto thy children, tibi hodie in corde tuo 6:7 et narrabis ea filiis tuis et medita-

and shalt talk of them when thou sittest in thine house, and when thou walkest by the way, and when thou liest down,

and when thou risest up. beris sedens in domo tua et ambulans in itinere dormiens at-6:8 And thou shalt bind them for a sign upon thine hand, que consurgens 6:8 et ligabis ea quasi signum in manu tua

and they shall be as frontlets between thine eyes.

6:9 And thou shalt write them upon the posts of thy house, eruntque et mouebuntur inter oculos tuos 6:9 scribesque ea

and on thy gates.

6:10 And it shall be, when the LORD thy God shall have in limine et ostiis domus tuae 6:10 cumque introduxerit te

brought thee into the land which he sware unto thy fathers, to Abraham, to Isaac, and to Jacob, to give thee great and

goodly cities, which thou buildedst not, Dominus Deus tuus in terram pro qua iurauit patribus tuis

Abraham Isaac et Iacob et dederit tibi ciuitates magnas et

6:11 And houses full of all good things, which thou filledst optimas quas non aedificasti 6:11 domos plenas cunctarum not, and wells digged, which thou diggedst not, vineyards and olive trees, which thou plantedst not; when thou shalt

have eaten and be full; opum quas non extruxisti cisternas quas non fodisti uineta 6:12 Then beware lest thou forget the LORD, which et oliueta quae non plantasti 6:12 et comederis et satura-

brought thee forth out of the land of Egypt, from the house of bondage.

tus fueris 6:13 caue diligenter ne obliuiscaris Domini qui

6:13 Thou shalt fear the LORD thy God, and serve him, and shalt swear by his name.

eduxit te de terra Aegypti de domo seruitutis Dominum Deum tuum timebis et ipsi seruies ac per nomen illius iurabis

6:14 Ye shall not go after other gods, of the gods of the 6:14 non ibitis post deos alienos cunctarum gentium quae in

people which are round about you;

6:15 (For the LORD thy God is a jealous God among you) circuitu uestro sunt 6:15 quoniam Deus aemulator Domi-

lest the anger of the LORD thy God be kindled against thee, and destroy thee from off the face of the earth.

nus Deus tuus in medio tui nequando irascatur furor Domini

6:16 Ye shall not tempt the LORD your God, as ye Dei tui contra te et auferat te de superficie terrae 6:16 non

tempted him in Massah.

temptabis Dominum Deum tuum sicut temptasti in loco tem-

6:17 Ye shall diligently keep the commandments of the ptationis 6:17 custodi praecepta Domini Dei tui ac testimo-

LORD your God, and his testimonies, and his statutes, which he hath commanded thee.

nia et caerimonias quas praecepit tibi 6:18 et fac quod pla-

6:18 And thou shalt do that which is right and good in the sight of the LORD: that it may be well with thee, and that thou mayest go in and possess the good land which the LORD sware unto thy fathers.

citum est et bonum in conspectu Domini ut bene sit tibi et ingressus possideas terram optimam de qua iurauit Dominus

6:19 To cast out all thine enemies from before thee, as the patribus tuis 6:19 ut deleret omnes inimicos tuos coram te

LORD hath spoken.

6:20 And when thy son asketh thee in time to come, say- sicut locutus est 6:20 cum interrogauerit te filius tuus cras

ing, What mean the testimonies, and the statutes, and the judgments, which the LORD our God hath commanded

you? dicens quid sibi uolunt testimonia haec et caerimoniae atque 6:21 Then thou shalt say unto thy son, We were Pharaoh’s iudicia quae praecepit Dominus Deus noster nobis 6:21 di-

bondmen in Egypt; and the LORD brought us out of Egypt with a mighty hand:

ces ei serui eramus Pharaonis in Aegypto et eduxit nos Do-

6:22 And the LORD shewed signs and wonders, great and minus de Aegypto in manu forti 6:22 fecitque signa atque

sore, upon Egypt, upon Pharaoh, and upon all his household, before our eyes:

Liber Deuteronomii

prodiiia magna et pessima in Aegypto contra Pharaonem et omnem domum illius in conspectu nostro 6:23 et eduxit nos 6:23 And he brought us out from thence, that he might

bring us in, to give us the land which he sware unto our fathers.

inde ut introductis daret terram super qua iurauit patribus nostris 6:24 praecepitque nobis Dominus ut faciamus omnia 6:24 And the LORD commanded us to do all these

statutes, to fear the LORD our God, for our good always, that he might preserve us alive, as it is at this day.

legitima haec et timeamus Dominum Deum nostrum et bene sit nobis cunctis diebus uitae nostrae sicut est hodie 6:25 6:25 And it shall be our righteousness, if we observe to do

all these commandments before the LORD our God, as he hath commanded us.

eritque nostri misericors si custodierimus et fecerimus omnia praecepta eius coram Domino Deo nostro sicut mandauit nobis

7:1 cum introduxerit te Dominus Deus tuus in terram 7:1 When the LORD thy God shall bring thee into the land whither thou goest to possess it, and hath cast out many nations before thee, the Hittites, and the Girgashites, and the Amorites, and the Canaanites, and the Perizzites, and the Hivites, and the Jebusites, seven nations greater and mightier than thou;

quam possessurus ingredieris et deleuerit gentes multas coram te Hettheum et Gergeseum et Amorreum Chananeum et Ferezeum et Eueum et Iebuseum septem gentes multo maioris numeri quam tu es et robustiores te 7:2 tradideritque eas 7:2 And when the LORD thy God shall deliver them be-

fore thee; thou shalt smite them, and utterly destroy them; thou shalt make no covenant with them, nor shew mercy

Dominus Deus tuus tibi percuties eas usque ad internicio- unto them: nem non inibis cum eis foedus nec misereberis earum 7:3 7:3 Neither shalt thou make marriages with them; thy

daughter thou shalt not give unto his son, nor his daughter shalt thou take unto thy son.

neque sociabis cum eis coniugia filiam tuam non dabis filio eius nec filiam illius accipies filio tuo 7:4 quia seducet 7:4 For they will turn away thy son from following me,

that they may serve other gods: so will the anger of the LORD be kindled against you, and destroy thee suddenly.

filium tuum ne sequatur me et ut magis seruiat diis alienis irasceturque furor Domini et delebit te cito 7:5 quin potius 7:5 But thus shall ye deal with them; ye shall destroy their

altars, and break down their images, and cut down their groves, and burn their graven images with fire.

haec facietis eis aras eorum subuertite confringite statuas lucosque succidite et sculptilia conburite 7:6 quia populus 7:6 For thou art an holy people unto the LORD thy God:

the LORD thy God hath chosen thee to be a special people unto himself, above all people that are upon the face of the

sanctus es Domino Deo tuo te elegit Dominus Deus tuus ut earth. sis ei populus peculiaris de cunctis populis qui sunt super terram 7:7 non quia cunctas gentes numero uincebatis uo- 7:7 The LORD did not set his love upon you, nor choose

you, because ye were more in number than any people; for ye were the fewest of all people:

bis iunctus est Dominus et elegit uos cum omnibus sitis populis pauciores 7:8 sed quia dilexit uos Dominus et custo- 7:8 But because the LORD loved you, and because he

would keep the oath which he had sworn unto your fathers, hath the LORD brought you out with a mighty hand, and redeemed you out of the house of bondmen, from the hand of Pharaoh king of Egypt.

diuit iuramentum quod iurauit patribus uestris eduxitque uos in manu forti et redeeit de domo seruitutis de manu Pharaonis regis Aegypti 7:9 et scies quia Dominus Deus tuus ipse 7:9 Know therefore that the LORD thy God, he is God,

the faithful God, which keepeth covenant and mercy with them that love him and keep his commandments to a thou-

est Deus fortis et fidelis custodiens pactum et misericordiam sand generations; diligentibus se et his qui custodiunt praecepta eius in mille generationes 7:10 et reddens odientibus se statim ita ut di- 7:10 And repayeth them that hate him to their face, to de-

stroy them: he will not be slack to him that hateth him, he will repay him to his face.

Biblia Sacra Vulgata

sperdat eos et ultra non differat protinus eis restituens quod

7:11 Thou shalt therefore keep the commandments, and merentur 7:11 custodi ergo praecepta et caerimonias atque

the statutes, and the judgments, which I command thee this day, to do them.

iudicia quae ego mando tibi hodie ut facias 7:12 si post-

7:12 Wherefore it shall come to pass, if ye hearken to these judgments, and keep, and do them, that the LORD thy God shall keep unto thee the covenant and the mercy which he sware unto thy fathers:

quam audieris haec iudicia custodieris ea et feceris custodiet et Dominus Deus tuus tibi pactum et misericordiam quam

7:13 And he will love thee, and bless thee, and multiply iurauit patribus tuis 7:13 et diliget te ac multiplicabit be-

thee: he will also bless the fruit of thy womb, and the fruit of thy land, thy corn, and thy wine, and thine oil, the increase of thy kine, and the flocks of thy sheep, in the land which he sware unto thy fathers to give thee.

nedicetque fructui uentris tui et fructui terrae tuae frumento tuo atque uindemiae oleo et armentis gregibus ouium tuarum super terram pro qua iurauit patribus tuis ut daret eam tibi

7:14 Thou shalt be blessed above all people: there shall 7:14 benedictus eris inter omnes populos non erit apud te ste-

not be male or female barren among you, or among your cattle.

rilis utriusque sexus tam in hominibus quam in gregibus tuis

7:15 And the LORD will take away from thee all sickness, 7:15 auferet Dominus a te omnem languorem et infirmitates

and will put none of the evil diseases of Egypt, which thou knowest, upon thee; but will lay them upon all them that

hate thee. Aegypti pessimas quas nouisti non inferet tibi sed cunctis 7:16 And thou shalt consume all the people which the hostibus tuis 7:16 deuorabis omnes populos quos Dominus

LORD thy God shall deliver thee; thine eye shall have no pity upon them: neither shalt thou serve their gods; for

that will be a snare unto thee. Deus tuus daturus est tibi non parcet eis oculus tuus nec ser-

7:17 If thou shalt say in thine heart, These nations are more uies diis eorum ne sint in ruinam tui 7:17 si dixeris in corde than I; how can I dispossess them?

tuo plures sunt gentes istae quam ego quomodo potero delere

7:18 Thou shalt not be afraid of them: but shalt well re- eas 7:18 noli metuere sed recordare quae fecerit Dominus

member what the LORD thy God did unto Pharaoh, and unto all Egypt;

Deus tuus Pharaoni et cunctis Aegyptiis 7:19 plagas maxi-

7:19 The great temptations which thine eyes saw, and the signs, and the wonders, and the mighty hand, and the stretched out arm, whereby the LORD thy God brought thee out: so shall the LORD thy God do unto all the people of whom thou art afraid.

mas quas uiderunt oculi tui et signa atque portenta manumque robustam et extentum brachium ut educeret te Dominus

7:20 Moreover the LORD thy God will send the hornet Deus tuus sic faciet cunctis populis quos metuis 7:20 insu-

among them, until they that are left, and hide themselves from thee, be destroyed.

per et crabrones mittet Dominus Deus tuus in eos donec deleat omnes atque disperdat qui te fugerint et latere potuerint

7:21 Thou shalt not be affrighted at them: for the LORD 7:21 non timebis eos quia Dominus Deus tuus in medio tui

thy God is among you, a mighty God and terrible.

7:22 And the LORD thy God will put out those nations est Deus magnus et terribilis 7:22 ipse consumet nationes

before thee by little and little: thou mayest not consume them at once, lest the beasts of the field increase upon thee.

has in conspectu tuo paulatim atque per partes non poteris delere eas pariter ne forte multiplicentur contra te bestiae

7:23 But the LORD thy God shall deliver them unto thee, terrae 7:23 dabitque eos Dominus Deus tuus in conspectu

and shall destroy them with a mighty destruction, until they be destroyed.

tuo et interficiet illos donec penitus deleantur 7:24 tradet

7:24 And he shall deliver their kings into thine hand, and thou shalt destroy their name from under heaven: there shall no man be able to stand before thee, until thou have destroyed them.

reges eorum in manus tuas et disperdes nomina eorum sub

7:25 The graven images of their gods shall ye burn with caelo nullus poterit resistere tibi donec conteras eos 7:25

fire: thou shalt not desire the silver or gold that is on them, nor take it unto thee, lest thou be snared therin: for it is an

abomination to the LORD thy God. sculptilia eorum igne conbures non concupisces argentum et

Liber Deuteronomii

aurum de quibus facta sunt neque adsumes ex eis tibi quicquam ne offendas propter ea quia abominatio est Domini Dei

tui 7:26 nec inferes quippiam ex idolo in domum tuam ne 7:26 Neither shalt thou bring an abomination into thine house, lest thou be a cursed thing like it: but thou shalt utterly detest it, and thou shalt utterly abhor it; for it is a

fias anathema sicut et illud est quasi spurcitiam detestaberis cursed thing. et uelut inquinamentum ac sordes abominationi habebis quia anathema est

8:1 omne mandatum quod ego praecipio tibi hodie caue di- 8:1 All the commandments which I command thee this day

shall ye observe to do, that ye may live, and multiply, and go in and possess the land which the LORD sware unto

ligenter ut facias ut possitis uiuere et multiplicemini ingres- your fathers. sique possideatis terram pro qua iurauit Dominus patribus uestris 8:2 et recordaberis cuncti itineris per quod adduxit 8:2 And thou shalt remember all the way which the LORD

thy God led thee these forty years in the wilderness, to humble thee, and to prove thee, to know what was in thine heart, whether thou wouldest keep his commandments, or no.

te Dominus Deus tuus quadraginta annis per desertum ut adfligeret te atque temptaret et nota fierent quae in tuo animo uersabantur utrum custodires mandata illius an non 8:3 ad- 8:3 And he humbled thee, and suffered thee to hunger,

and fed thee with manna, which thou knewest not, neither did thy fathers know; that he might make thee know that man doth not live by bread only, but by every word that proceedeth out of the mouth of the LORD doth man live.

flixit te penuria et dedit tibi cibum manna quem ignorabas tu et patres tui ut ostenderet tibi quod non in solo pane uiuat

homo sed in omni uerbo quod egreditur ex ore Domini 8:4 8:4 Thy raiment waxed not old upon thee, neither did thy

foot swell, these forty years.

uestimentum tuum quo operiebaris nequaquam uetustate defecit et pes tuus non est subtritus en quadragesimus annus est 8:5 ut recogites in corde tuo quia sicut erudit homo filium 8:5 Thou shalt also consider in thine heart, that, as a man

chasteneth his son, so the LORD thy God chasteneth thee.

suum sic Dominus Deus tuus erudiuit te 8:6 ut custodias 8:6 Therefore thou shalt keep the commandments of the

LORD thy God, to walk in his ways, and to fear him.

mandata Domini Dei tui et ambules in uiis eius et timeas eum 8:7 Dominus enim Deus tuus introducet te in terram bonam 8:7 For the LORD thy God bringeth thee into a good land,

a land of brooks of water, of fountains and depths that spring out of valleys and hills;

terram riuorum aquarumque et fontium in cuius campis et montibus erumpunt fluuiorum abyssi 8:8 terram frumenti 8:8 A land of wheat, and barley, and vines, and fig trees,

and pomegranates; a land of oil olive, and honey;

hordei uinearum in qua ficus et mala granata et oliueta nascuntur terram olei ac mellis 8:9 ubi absque ulla penuria co- 8:9 A land wherein thou shalt eat bread without scarce-

ness, thou shalt not lack any thing in it; a land whose stones are iron, and out of whose hills thou mayest dig

medes panem tuum et rerum omnium abundantia perfrueris brass. cuius lapides ferrum sunt et de montibus eius aeris metalla fodiuntur 8:10 ut cum comederis et satiatus fueris benedi- 8:10 When thou hast eaten and art full, then thou shalt

bless the LORD thy God for the good land which he hath given thee.

cas Domino Deo tuo pro terra optima quam dedit tibi 8:11

8:11 Beware that thou forget not the LORD thy God, in not keeping his commandments, and his judgments, and his statutes, which I command thee this day:

obserua et caue nequando obliuiscaris Domini Dei tui et neglegas mandata eius atque iudicia et caerimonias quas ego praecipio tibi hodie 8:12 ne postquam comederis et satia- 8:12 Lest when thou hast eaten and art full, and hast built

goodly houses, and dwelt therein;

Biblia Sacra Vulgata

8:13 And when thy herds and thy flocks multiply, and thy tus domos pulchras aedificaueris et habitaueris in eis 8:13

silver and thy gold is multiplied, and all that thou hast is multiplied;

habuerisque armenta et ouium greges argenti et auri cunc-

8:14 Then thine heart be lifted up, and thou forget the tarumque rerum copiam 8:14 eleuetur cor tuum et non re-

LORD thy God, which brought thee forth out of the land of Egypt, from the house of bondage;

miniscaris Domini Dei tui qui eduxit te de terra Aegypti de

8:15 Who led thee through that great and terrible wilder- domo seruitutis 8:15 et ductor tuus fuit in solitudine magna

ness, wherein were fiery serpents, and scorpions, and drought, where there was no water; who brought thee forth

water out of the rock of flint; atque terribili in qua erat serpens flatu adurens et scorpio ac

dipsas et nullae omnino aquae qui eduxit riuos de petra du-

8:16 Who fed thee in the wilderness with manna, which rissima 8:16 et cibauit te manna in solitudine quod nescie-

thy fathers knew not, that he might humble thee, and that he might prove thee, to do thee good at thy latter end;

runt patres tui et postquam adflixit ac probauit ad extremum

8:17 And thou say in thine heart, My power and the might misertus est tui 8:17 ne diceres in corde tuo fortitudo mea

of mine hand hath gotten me this wealth.

8:18 But thou shalt remember the LORD thy God: for it et robur manus meae haec mihi omnia praestiterunt 8:18

is he that giveth thee power to get wealth, that he may establish his covenant which he sware unto thy fathers, as

it is this day. sed recorderis Domini Dei tui quod ipse tibi uires praebuerit

ut impleret pactum suum super quo iurauit patribus tuis sic-

8:19 And it shall be, if thou do at all forget the LORD ut praesens indicat dies 8:19 sin autem oblitus Domini Dei

thy God, and walk after other gods, and serve them, and worship them, I testify against you this day that ye shall

surely perish. tui secutus fueris alienos deos coluerisque illos et adoraueris 8:20 As the nations which the LORD destroyeth before ecce nunc praedico tibi quod omnino dispereas 8:20 sicut

your face, so shall ye perish; because ye would not be obedient unto the voice of the LORD your God.

gentes quas deleuit Dominus in introitu tuo ita et uos peribitis si inoboedientes fueritis uoci Domini Dei uestri

9:1 Hear, O Israel: Thou art to pass over Jordan this day, to 9:1 audi Israhel tu transgredieris hodie Iordanem ut posgo in to possess nations greater and mightier than thyself, cities great and fenced up to heaven,

sideas nationes maximas et fortiores te ciuitates ingentes et

9:2 A people great and tall, the children of the Anakims, ad caelum usque muratas 9:2 populum magnum atque suwhom thou knowest, and of whom thou hast heard say, Who can stand before the children of Anak!

blimem filios Enacim quos ipse uidisti et audisti quibus nul-

9:3 Understand therefore this day, that the LORD thy God lus potest ex aduerso resistere 9:3 scies ergo hodie quod

is he which goeth over before thee; as a consuming fire he shall destroy them, and he shall bring them down before thy face: so shalt thou drive them out, and destroy them quickly, as the LORD hath said unto thee.

Dominus Deus tuus ipse transibit ante te ignis deuorans atque consumens qui conterat eos et deleat atque disperdat

9:4 Speak not thou in thine heart, after that the LORD ante faciem tuam uelociter sicut locutus est tibi 9:4 ne di-

thy God hath cast them out from before thee, saying, For my righteousness the LORD hath brought me in to possess this land: but for the wickedness of these nations the LORD doth drive them out from before thee.

cas in corde tuo cum deleuerit eos Dominus Deus tuus in conspectu tuo propter iustitiam meam introduxit me Dominus ut terram hanc possiderem cum propter impietates suas

9:5 Not for thy righteousness, or for the uprightness of istae deletae sint nationes 9:5 neque enim propter iustitias

thine heart, dost thou go to possess their land: but for the wickedness of these nations the LORD thy God doth drive them out from before thee, and that he may perform the word which the LORD sware unto thy fathers, Abraham, Isaac, and Jacob.

tuas et aequitatem cordis tui ingredieris ut possideas terras eorum sed quia illae egerunt impie te introeunte deletae sunt et ut conpleret uerbum suum Dominus quod sub iuramento

Liber Deuteronomii

pollicitus est patribus tuis Abraham Isaac et Iacob 9:6 sc- 9:6 Understand therefore, that the LORD thy God giveth

thee not this good land to possess it for thy righteousness; for thou art a stiffnecked people.

ito igitur quod non propter iustitias tuas Dominus Deus tuus dederit tibi terram hanc optimam in possessionem cum durissimae ceruicis sis populus 9:7 memento et ne obliuisca- 9:7 Remember, and forget not, how thou provokedst the

LORD thy God to wrath in the wilderness: from the day that thou didst depart out of the land of Egypt, until ye came unto this place, ye have been rebellious against the LORD.

ris quomodo ad iracundiam prouocaueris Dominum Deum tuum in solitudine ex eo die quo es egressus ex Aegypto usque ad locum istum semper aduersum Dominum contendisti 9:8 nam et in Horeb prouocasti eum et iratus delere te uoluit 9:8 Also in Horeb ye provoked the LORD to wrath, so that

the LORD was angry with you to have destroyed you.

9:9 quando ascendi in montem ut acciperem tabulas lapideas 9:9 When I was gone up into the mount to receive the ta-

bles of stone, even the tables of the covenant which the LORD made with you, then I abode in the mount forty days and forty nights, I neither did eat bread nor drink water:

tabulas pacti quod pepigit uobiscum Dominus et perseueraui in monte quadraginta diebus ac noctibus panem non comedens et aquam non bibens 9:10 deditque mihi Dominus 9:10 And the LORD delivered unto me two tables of stone

written with the finger of God; and on them was written according to all the words, which the LORD spake with you in the mount out of the midst of the fire in the day of the assembly.

duas tabulas lapideas scriptas digito Dei et continentes omnia uerba quae uobis in monte locutus est de medio ignis quando contio populi congregata est 9:11 cumque transis- 9:11 And it came to pass at the end of forty days and forty

nights, that the LORD gave me the two tables of stone, even the tables of the covenant.

sent quadraginta dies et totidem noctes dedit mihi Dominus duas tabulas lapideas tabulas foederis 9:12 dixitque mihi 9:12 And the LORD said unto me, Arise, get thee down

quickly from hence; for thy people which thou hast brought forth out of Egypt have corrupted themselves; they are quickly turned aside out of the way which I commanded them; they have made them a molten image.

surge et descende hinc cito quia populus tuus quos eduxisti de Aegypto deseruerunt uelociter uiam quam demonstrasti eis feceruntque sibi conflatile 9:13 rursumque ait Dominus 9:13 Furthermore the LORD spake unto me, saying, I have

seen this people, and, behold, it is a stiffnecked people:

ad me cerno quod populus iste durae ceruicis sit 9:14 di- 9:14 Let me alone, that I may destroy them, and blot out

their name from under heaven: and I will make of thee a nation mightier and greater than they.

mitte me ut conteram eum et deleam nomen eius sub caelo et constituam te super gentem quae hac maior et fortior sit 9:15 cumque de monte ardente descenderem et duas tabulas 9:15 So I turned and came down from the mount, and the

mount burned with fire: and the two tables of the covenant were in my two hands.

foederis utraque tenerem manu 9:16 uidissemque uos pec-

9:16 And I looked, and, behold, ye had sinned against the LORD your God, and had made you a molten calf: ye had turned aside quickly out of the way which the LORD had commanded you.

casse Domino Deo uestro et fecisse uobis uitulum conflatilem ac deseruisse uelociter uiam eius quam uobis ostenderat 9:17 proieci tabulas de manibus meis confregique eas in con- 9:17 And I took the two tables, and cast them out of my

two hands, and brake them before your eyes.

spectu uestro 9:18 et procidi ante Dominum sicut prius qua- 9:18 And I fell down before the LORD, as at the first, forty

days and forty nights: I did neither eat bread, nor drink water, because of all your sins which ye sinned, in doing wickedly in the sight of the LORD, to provoke him to anger.

draginta diebus et noctibus panem non comedens et aquam non bibens propter omnia peccata uestra quae gessistis contra Dominum et eum ad iracundiam prouocastis 9:19 timui 9:19 For I was afraid of the anger and hot displeasure,

wherewith the LORD was wroth against you to destroy you. But the LORD hearkened unto me at that time also.

enim indignationem et iram illius qua aduersum uos conci-

Biblia Sacra Vulgata

tatus delere uos uoluit et exaudiuit me Dominus etiam hac

9:20 And the LORD was very angry with Aaron to have uice 9:20 aduersum Aaron quoque uehementer iratus uo-

destroyed him: and I prayed for Aaron also the same time.

9:21 And I took your sin, the calf which ye had made, luit eum conterere et pro illo similiter deprecatus sum 9:21

and burnt it with fire, and stamped it, and ground it very small, even until it was as small as dust: and I cast the dust

thereof into the brook that descended out of the mount. peccatum autem uestrum quod feceratis id est uitulum arripiens igne conbusi et in frusta comminuens omninoque in puluerem rediiens proieci in torrentem qui de monte descen-

9:22 And at Taberah, and at Massah, and at Kibrothhat- dit 9:22 in Incendio quoque et in Temptatione et in sepulch-

taavah, ye provoked the LORD to wrath.

9:23 Likewise when the LORD sent you from Kadesh- ris Concupiscentiae prouocastis Dominum 9:23 et quando

barnea, saying, Go up and possess the land which I have given you; then ye rebelled against the commandment of the LORD your God, and ye believed him not, nor hearkened to his voice.

misit uos de Cadesbarne dicens ascendite et possidete terram quam dedi uobis et contempsistis imperium Domini Dei uestri et non credidistis ei neque uocem eius audire uoluistis

9:24 Ye have been rebellious against the LORD from the 9:24 sed semper fuistis rebelles a die qua nosse uos coepi

day that I knew you.

9:25 Thus I fell down before the LORD forty days and 9:25 et iacui coram Domino quadraginta diebus ac noctibus

forty nights, as I fell down at the first; because the LORD had said he would destroy you.

quibus eum suppliciter deprecabar ne deleret uos ut fuerat

9:26 I prayed therefore unto the LORD, and said, O Lord comminatus 9:26 et orans dixi Domine Deus ne disperdas

GOD, destroy not thy people and thine inheritance, which thou hast redeemed through thy greatness, which thou hast

brought forth out of Egypt with a mighty hand. populum tuum et hereditatem tuam quam redeeisti in mag-

9:27 Remember thy servants, Abraham, Isaac, and Jacob; nitudine tua quos eduxisti de Aegypto in manu forti 9:27 look not unto the stubbornness of this people, nor to their wickedness, nor to their sin:

recordare seruorum tuorum Abraham Isaac et Iacob ne aspicias duritiam populi huius et impietatem atque peccatum

9:28 Lest the land whence thou broughtest us out say, Be- 9:28 ne forte dicant habitatores terrae de qua eduxisti nos

cause the LORD was not able to bring them into the land which he promised them, and because he hated them, he

hath brought them out to slay them in the wilderness. non poterat Dominus introducere eos in terram quam polli-

citus est eis et oderat illos idcirco eduxit ut interficeret eos in

9:29 Yet they are thy people and thine inheritance, which solitudine 9:29 qui sunt populus tuus et hereditas tua quos

thou broughtest out by thy mighty power and by thy stretched out arm.

eduxisti in fortitudine tua magna et in brachio tuo extento

10:1 At that time the LORD said unto me, Hew thee two 10:1 in tempore illo dixit Dominus ad me dola tibi duas

tables of stone like unto the first, and come up unto me into the mount, and make thee an ark of wood.

tabulas lapideas sicut priores fuerunt et ascende ad me in

10:2 And I will write on the tables the words that were in montem faciesque arcam ligneam 10:2 et scribam in tabu-

the first tables which thou brakest, and thou shalt put them in the ark.

lis uerba quae fuerunt in his quas ante confregisti ponesque

10:3 And I made an ark of shittim wood, and hewed two eas in arca 10:3 feci igitur arcam de lignis setthim cum-

tables of stone like unto the first, and went up into the mount, having the two tables in mine hand.

que dolassem duas tabulas lapideas instar priorum ascendi

10:4 And he wrote on the tables, according to the first writ- in montem habens eas in manibus 10:4 scripsitque in tabuing, the ten commandments, which the LORD spake unto you in the mount out of the midst of the fire in the day of

the assembly: and the LORD gave them unto me. lis iuxta id quod prius scripserat uerba decem quae locutus est Dominus ad uos in monte de medio ignis quando po-

Liber Deuteronomii

pulus congregatus est et dedit eas mihi 10:5 reuersusque 10:5 And I turned myself and came down from the mount,

and put the tables in the ark which I had made; and there they be, as the LORD commanded me.

de monte descendi et posui tabulas in arcam quam feceram quae hucusque ibi sunt sicut mihi praecepit Dominus 10:6 10:6 And the children of Israel took their journey from

Beeroth of the children of Jaakan to Mosera: there Aaron died, and there he was buried; and Eleazar his son minis-

filii autem Israhel castra mouerunt ex Beroth filiorum Iacan tered in the priest’s office in his stead. in Musera ubi Aaron mortuus ac sepultus est pro quo sacerdotio functus est filius eius Eleazar 10:7 inde uenerunt in 10:7 From thence they journeyed unto Gudgodah; and

from Gudgodah to Jotbath, a land of rivers of waters.

Gadgad de quo loco profecti castrametati sunt in Ietabatha in terra aquarum atque torrentium 10:8 eo tempore separa- 10:8 At that time the LORD separated the tribe of Levi, to

bear the ark of the covenant of the LORD, to stand before the LORD to minister unto him, and to bless in his name,

uit tribum Leui ut portaret arcam foederis Domini et staret unto this day. coram eo in ministerio ac benediceret in nomine illius usque

in praesentem diem 10:9 quam ob rem non habuit Leui par- 10:9 Wherefore Levi hath no part nor inheritance with his brethren; the LORD is his inheritance, according as the LORD thy God promised him.

tem neque possessionem cum fratribus suis quia ipse Dominus possessio eius est sicut promisit ei Dominus Deus tuus 10:10 ego autem steti in monte sicut prius quadraginta diebus 10:10 And I stayed in the mount, according to the first

time, forty days and forty nights; and the LORD hearkened unto me at that time also, and the LORD would not

ac noctibus exaudiuitque me Dominus etiam hac uice et te destroy thee. perdere noluit 10:11 dixitque mihi uade et praecede popu- 10:11 And the LORD said unto me, Arise, take thy jour-

ney before the people, that they may go in and possess the land, which I sware unto their fathers to give unto them.

lum ut ingrediatur et possideat terram quam iuraui patribus eorum ut traderem eis 10:12 et nunc Israhel quid Dominus 10:12 And now, Israel, what doth the LORD thy God re-

quire of thee, but to fear the LORD thy God, to walk in all his ways, and to love him, and to serve the LORD thy God

Deus tuus petit a te nisi ut timeas Dominum Deum tuum et with all thy heart and with all thy soul, ambules in uiis eius et diligas eum ac seruias Domino Deo tuo in toto corde tuo et in tota anima tua 10:13 custodiasque 10:13 To keep the commandments of the LORD, and his

statutes, which I command thee this day for thy good?

mandata Domini et caerimonias eius quas ego hodie praeci-

pio ut bene sit tibi 10:14 en Domini Dei tui caelum est et 10:14 Behold, the heaven and the heaven of heavens is the

LORD’s thy God, the earth also, with all that therein is.

caelum caeli terra et omnia quae in ea sunt 10:15 et tamen 10:15 Only the LORD had a delight in thy fathers to love

them, and he chose their seed after them, even you above all people, as it is this day.

patribus tuis conglutinatus est Dominus et amauit eos elegitque semen eorum post eos id est uos de cunctis gentibus sicut hodie conprobatur 10:16 circumcidite igitur praepu- 10:16 Circumcise therefore the foreskin of your heart, and

be no more stiffnecked.

tium cordis uestri et ceruicem uestram ne induretis amplius 10:17 quia Dominus Deus uester ipse est Deus deorum et Do- 10:17 For the LORD your God is God of gods, and Lord

of lords, a great God, a mighty, and a terrible, which regardeth not persons, nor taketh reward:

minus dominantium Deus magnus et potens et terribilis qui personam non accipit nec munera 10:18 facit iudicium pu- 10:18 He doth execute the judgment of the fatherless and

widow, and loveth the stranger, in giving him food and raiment.

pillo et uiduae amat peregrinum et dat ei uictum atque uestitum 10:19 et uos ergo amate peregrinos quia et ipsi fuistis 10:19 Love ye therefore the stranger: for ye were strangers

in the land of Egypt.

Biblia Sacra Vulgata

10:20 Thou shalt fear the LORD thy God; him shalt thou aduenae in terra Aegypti 10:20 Dominum Deum tuum tiserve, and to him shalt thou cleave, and swear by his name.

mebis et ei seruies ipsi adherebis iurabisque in nomine illius

10:21 He is thy praise, and he is thy God, that hath done 10:21 ipse est laus tua et Deus tuus qui fecit tibi haec magna-

for thee these great and terrible things, which thine eyes have seen.

lia et terribilia quae uiderunt oculi tui 10:22 in septuaginta

10:22 Thy fathers went down into Egypt with threescore and ten persons; and now the LORD thy God hath made thee as the stars of heaven for multitude.

animabus descenderunt patres tui in Aegyptum et ecce nunc multiplicauit te Dominus Deus tuus sicut astra caeli

11:1 Therefore thou shalt love the LORD thy God, and 11:1 ama itaque Dominum Deum tuum et obserua praekeep his charge, and his statutes, and his judgments, and his commandments, alway.

cepta eius et caerimonias iudicia atque mandata omni tem-

11:2 And know ye this day: for I speak not with your chil- pore 11:2 cognoscite hodie quae ignorant filii uestri qui non

dren which have not known, and which have not seen the chastisement of the LORD your God, his greatness, his

mighty hand, and his stretched out arm, uiderunt disciplinam Domini Dei uestri magnalia eius et ro-11:3 And his miracles, and his acts, which he did in the bustam manum extentumque brachium 11:3 signa et opera

midst of Egypt unto Pharaoh the king of Egypt, and unto all his land;

quae fecit in medio Aegypti Pharaoni regi et uniuersae terrae

11:4 And what he did unto the army of Egypt, unto their eius 11:4 omnique exercitui Aegyptiorum et equis ac cur-

horses, and to their chariots; how he made the water of the Red sea to overflow them as they pursued after you, and

how the LORD hath destroyed them unto this day; ribus quomodo operuerint eos aquae Rubri maris cum uos

persequerentur et deleuerit eos Dominus usque in praesen-

11:5 And what he did unto you in the wilderness, until ye tem diem 11:5 uobisque quae fecerit in solitudine donec ue-

came into this place;

11:6 And what he did unto Dathan and Abiram, the sons of niretis ad hunc locum 11:6 et Dathan atque Abiram filiis

Eliab, the son of Reuben: how the earth opened her mouth, and swallowed them up, and their households, and their tents, and all the substance that was in their possession, in the midst of all Israel:

Heliab qui fuit filius Ruben quos aperto ore suo terra absorbuit cum domibus et tabernaculis et uniuersa substantia

11:7 But your eyes have seen all the great acts of the eorum quam habebant in medio Israhelis 11:7 oculi uestri

LORD which he did. 11:8 Therefore shall ye keep all the commandments which uiderunt omnia opera Domini magna quae fecit 11:8 ut cu-I command you this day, that ye may be strong, and go in and possess the land, whither ye go to possess it;

stodiatis uniuersa mandata illius quae ego hodie praecipio uobis et possitis introire et possidere terram ad quam ingre-

11:9 And that ye may prolong your days in the land, which dimini 11:9 multoque in ea uiuatis tempore quam sub iura-

the LORD sware unto your fathers to give unto them and to their seed, a land that floweth with milk and honey.

mento pollicitus est Dominus patribus uestris et semini eo-

11:10 For the land, whither thou goest in to possess it, is rum lacte et melle manantem 11:10 terra enim ad quam innot as the land of Egypt, from whence ye came out, where thou sowedst thy seed, and wateredst it with thy foot, as a

garden of herbs: gredieris possidendam non est sicut terra Aegypti de qua ex-

isti ubi iacto semine in hortorum morem aquae ducuntur inri-

11:11 But the land, whither ye go to possess it, is a land of guae 11:11 sed montuosa est et campestris de caelo expechills and valleys, and drinketh water of the rain of heaven:

11:12 A land which the LORD thy God careth for: the tans pluuias 11:12 quam Dominus Deus tuus semper inuisit

eyes of the LORD thy God are always upon it, from the beginning of the year even unto the end of the year.

et oculi illius in ea sunt a principio anni usque ad finem eius

11:13 And it shall come to pass, if ye shall hearken dili- 11:13 si ergo oboedieritis mandatis meis quae hodie praeci-

gently unto my commandments which I command you this day, to love the LORD your God, and to serve him with all

your heart and with all your soul, pio uobis ut diligatis Dominum Deum uestrum et seruiatis

Liber Deuteronomii

ei in toto corde uestro et in tota anima uestra 11:14 dabo 11:14 That I will give you the rain of your land in his due

season, the first rain and the latter rain, that thou mayest gather in thy corn, and thy wine, and thine oil.

pluuiam terrae uestrae temporiuam et serotinam ut colliga-

tis frumentum et uinum et oleum 11:15 faenum ex agris ad 11:15 And I will send grass in thy fields for thy cattle, that

thou mayest eat and be full.

pascenda iumenta et ut ipsi comedatis ac saturemini 11:16 11:16 Take heed to yourselves, that your heart be not de-

ceived, and ye turn aside, and serve other gods, and worship them;

cauete ne forte decipiatur cor uestrum et recedatis a Domino seruiatisque diis alienis et adoretis eos 11:17 iratusque Do- 11:17 And then the LORD’s wrath be kindled against you,

and he shut up the heaven, that there be no rain, and that the land yield not her fruit; and lest ye perish quickly from

minus claudat caelum et pluuiae non descendant nec terra off the good land which the LORD giveth you. det germen suum pereatisque uelociter de terra optima quam Dominus daturus est uobis 11:18 ponite haec uerba mea in 11:18 Therefore shall ye lay up these my words in your

heart and in your soul, and bind them for a sign upon your hand, that they may be as frontlets between your eyes.

cordibus et in animis uestris et suspendite ea pro signo in manibus et inter uestros oculos conlocate 11:19 docete fi- 11:19 And ye shall teach them your children, speaking

of them when thou sittest in thine house, and when thou walkest by the way, when thou liest down, and when thou

lios uestros ut illa meditentur quando sederis in domo tua risest up. et ambulaueris in uia et accubueris atque surrexeris 11:20 11:20 And thou shalt write them upon the door posts of

thine house, and upon thy gates:

scribes ea super postes et ianuas domus tuae 11:21 ut mul- 11:21 That your days may be multiplied, and the days of

your children, in the land which the LORD sware unto your fathers to give them, as the days of heaven upon the

tiplicentur dies tui et filiorum tuorum in terra quam iurauit earth. Dominus patribus tuis ut daret eis quamdiu caelum inminet terrae 11:22 si enim custodieritis mandata quae ego praeci- 11:22 For if ye shall diligently keep all these command-

ments which I command you, to do them, to love the LORD your God, to walk in all his ways, and to cleave

pio uobis et feceritis ea ut diligatis Dominum Deum uestrum unto him; et ambuletis in omnibus uiis eius adherentes ei 11:23 dis- 11:23 Then will the LORD drive out all these nations from

before you, and ye shall possess greater nations and mightier than yourselves.

perdet Dominus omnes gentes istas ante faciem uestram et possidebitis eas quae maiores et fortiores uobis sunt 11:24 11:24 Every place whereon the soles of your feet shall

tread shall be yours: from the wilderness and Lebanon, from the river, the river Euphrates, even unto the uttermost

omnis locus quem calcauerit pes uester uester erit a deserto sea shall your coast be. et Libano a flumine magno Eufraten usque ad mare occidentale erunt termini uestri 11:25 nullus stabit contra uos 11:25 There shall no man be able to stand before you: for

the LORD your God shall lay the fear of you and the dread of you upon all the land that ye shall tread upon, as he hath

terrorem uestrum et formidinem dabit Dominus Deus ues- said unto you. ter super omnem terram quam calcaturi estis sicut locutus est uobis 11:26 en propono in conspectu uestro hodie bene- 11:26 Behold, I set before you this day a blessing and a

curse;

dictionem et maledictionem 11:27 benedictionem si oboe- 11:27 A blessing, if ye obey the commandments of the

LORD your God, which I command you this day:

dieritis mandatis Domini Dei uestri quae ego praecipio uobis 11:28 maledictionem si non audieritis mandata Domini 11:28 And a curse, if ye will not obey the commandments

of the LORD your God, but turn aside out of the way which I command you this day, to go after other gods,

Dei uestri sed recesseritis de uia quam ego nunc ostendo uo- which ye have not known. bis et ambulaueritis post deos alienos quos ignoratis 11:29 11:29 And it shall come to pass, when the LORD thy God

hath brought thee in unto the land whither thou goest to possess it, that thou shalt put the blessing upon mount Ger-

cum introduxerit te Dominus Deus tuus in terram ad quam izim, and the curse upon mount Ebal.

Biblia Sacra Vulgata

pergis habitandam pones benedictionem super montem Ga-

11:30 Are they not on the other side Jordan, by the way rizim maledictionem super montem Hebal 11:30 qui sunt

where the sun goeth down, in the land of the Canaanites, which dwell in the champaign over against Gilgal, beside

the plains of Moreh? trans Iordanem post uiam quae uergit ad solis occubitum in

terra Chananei qui habitat in campestribus contra Galgalam

11:31 For ye shall pass over Jordan to go in to possess the quae est iuxta uallem tendentem et intrantem procul 11:31

land which the LORD your God giveth you, and ye shall possess it, and dwell therein.

uos enim transibitis Iordanem ut possideatis terram quam Dominus Deus uester daturus est uobis et habeatis ac possi-

11:32 And ye shall observe to do all the statutes and judg- deatis illam 11:32 uidete ergo ut impleatis caerimonias at-

ments which I set before you this day.

que iudicia quae ego hodie ponam in conspectu uestro

12:1 These are the statutes and judgments, which ye shall 12:1 haec sunt praecepta atque iudicia quae facere debeobserve to do in the land, which the LORD God of thy fathers giveth thee to possess it, all the days that ye live

upon the earth. tis in terra quam Dominus Deus patrum tuorum daturus est

tibi ut possideas eam cunctis diebus quibus super humum

12:2 Ye shall utterly destroy all the places, wherein the gradieris 12:2 subuertite omnia loca in quibus coluerunt

nations which ye shall possess served their gods, upon the high mountains, and upon the hills, and under every green

tree: gentes quas possessuri estis deos suos super montes excel-

12:3 And ye shall overthrow their altars, and break their sos et colles et subter omne lignum frondosum 12:3 dissipillars, and burn their groves with fire; and ye shall hew down the graven images of their gods, and destroy the

names of them out of that place. pate aras earum et confringite statuas lucos igne conburite

et idola comminuite disperdite nomina eorum de locis illis

12:4 Ye shall not do so unto the LORD your God. 12:4 non facietis ita Domino Deo uestro 12:5 sed ad locum

12:5 But unto the place which the LORD your God shall choose out of all your tribes to put his name there, even unto his habitation shall ye seek, and thither thou shalt come:

quem elegerit Dominus Deus uester de cunctis tribubus ue-

12:6 And thither ye shall bring your burnt offerings, and stris ut ponat nomen suum ibi et habitet in eo uenietis 12:6

your sacrifices, and your tithes, and heave offerings of your hand, and your vows, and your freewill offerings, and

the firstlings of your herds and of your flocks: et offeretis in illo loco holocausta et uictimas uestras deci-

mas et primitias manuum uestrarum et uota atque donaria

12:7 And there ye shall eat before the LORD your God, primogenita boum et ouium 12:7 et comedetis ibi in con-

and ye shall rejoice in all that ye put your hand unto, ye and your households, wherein the LORD thy God hath

blessed thee. spectu Domini Dei uestri ac laetabimini in cunctis ad quae

miseritis manum uos et domus uestrae in quibus benedixerit

12:8 Ye shall not do after all the things that we do here this uobis Dominus Deus uester 12:8 non facietis ibi quae nos day, every man whatsoever is right in his own eyes.

12:9 For ye are not as yet come to the rest and to the in- hic facimus hodie singuli quod sibi rectum uidetur 12:9 ne-

heritance, which the LORD your God giveth you.

que enim usque in praesens tempus uenistis ad requiem et

12:10 But when ye go over Jordan, and dwell in the land possessionem quam Dominus Deus daturus est uobis 12:10

which the LORD your God giveth you to inherit, and when he giveth you rest from all your enemies round about, so

that ye dwell in safety; transibitis Iordanem et habitabitis in terram quam Dominus

Deus uester daturus est uobis ut requiescatis a cunctis hos-

12:11 Then there shall be a place which the LORD your tibus per circuitum et absque ullo timore habitetis 12:11 in

God shall choose to cause his name to dwell there; thither shall ye bring all that I command you; your burnt offerings, and your sacrifices, your tithes, and the heave offering of your hand, and all your choice vows which ye vow unto the LORD:

loco quem elegerit Dominus Deus uester ut sit nomen eius

Liber Deuteronomii

in eo illuc omnia quae praecipio conferetis holocausta et hostias ac decimas et primitias manuum uestrarum et quicquid praecipuum est in muneribus quae uouistis Domino 12:12 12:12 And ye shall rejoice before the LORD your God, ye,

and your sons, and your daughters, and your menservants, and your maidservants, and the Levite that is within your gates; forasmuch as he hath no part nor inheritance with you.

ibi epulabimini coram Domino Deo uestro uos filii ac filiae uestrae famuli et famulae atque Leuites qui in uestris urbibus commorantur neque enim habet aliam partem et possessionem inter uos 12:13 caue ne offeras holocausta tua in 12:13 Take heed to thyself that thou offer not thy burnt

offerings in every place that thou seest:

omni loco quem uideris 12:14 sed in eo quem elegerit Do- 12:14 But in the place which the LORD shall choose in

one of thy tribes, there thou shalt offer thy burnt offerings, and there thou shalt do all that I command thee.

minus in una tribuum tuarum offeres hostias et facies quaecumque praecipio tibi 12:15 sin autem comedere uolueris 12:15 Notwithstanding thou mayest kill and eat flesh in

all thy gates, whatsoever thy soul lusteth after, according to the blessing of the LORD thy God which he hath given thee: the unclean and the clean may eat thereof, as of the roebuck, and as of the hart.

et te esus carnium delectarit occide et comede iuxta benedictionem Domini Dei tui quam dedit tibi in urbibus tuis siue inmundum fuerit hoc est maculatum et debile siue mundum hoc est integrum et sine macula quod offerri licet sic-

ut capream et ceruum comedes 12:16 absque esu dumtaxat 12:16 Only ye shall not eat the blood; ye shall pour it upon

the earth as water.

sanguinis quod super terram quasi aquam effundes 12:17 12:17 Thou mayest not eat within thy gates the tithe of

thy corn, or of thy wine, or of thy oil, or the firstlings of thy herds or of thy flock, nor any of thy vows which thou vowest, nor thy freewill offerings, or heave offering of thine hand:

non poteris comedere in oppidis tuis decimam frumenti et uini et olei tui primogenita armentorum et pecorum et omnia quae uoueris et sponte offerre uolueris et primitias manuum tuarum 12:18 sed coram Domino Deo tuo comedes ea in 12:18 But thou must eat them before the LORD thy God

in the place which the LORD thy God shall choose, thou, and thy son, and thy daughter, and thy manservant, and thy maidservant, and the Levite that is within thy gates: and thou shalt rejoice before the LORD thy God in all that thou puttest thine hands unto.

loco quem elegerit Dominus Deus tuus tu et filius tuus ac filia seruus et famula atque Leuites qui manet in urbibus tuis et laetaberis et reficieris coram Domino Deo tuo in cunctis

ad quae extenderis manum tuam 12:19 caue ne derelinquas 12:19 Take heed to thyself that thou forsake not the Levite

as long as thou livest upon the earth.

Leuiten omni tempore quo uersaris in terra 12:20 quando 12:20 When the LORD thy God shall enlarge thy border,

as he hath promised thee, and thou shalt say, I will eat flesh, because thy soul longeth to eat flesh; thou mayest

dilatauerit Dominus Deus tuus terminos tuos sicut locutus eat flesh, whatsoever thy soul lusteth after. est tibi et uolueris uesci carnibus quas desiderat anima tua 12:21 locus autem quem elegerit Dominus Deus tuus ut sit 12:21 If the place which the LORD thy God hath chosen

to put his name there be too far from thee, then thou shalt kill of thy herd and of thy flock, which the LORD hath given thee, as I have commanded thee, and thou shalt eat in thy gates whatsoever thy soul lusteth after.

nomen eius ibi si procul fuerit occides de armentis et pecoribus quae habueris sicut praecepi tibi et comedes in oppidis tuis ut tibi placet 12:22 sicut comeditur caprea et ceruus 12:22 Even as the roebuck and the hart is eaten, so thou

shalt eat them: the unclean and the clean shall eat of them alike.

ita uesceris eis et mundus et inmundus in commune uescentur 12:23 hoc solum caue ne sanguinem comedas sanguis 12:23 Only be sure that thou eat not the blood: for the

blood is the life; and thou mayest not eat the life with the flesh.

Biblia Sacra Vulgata

enim eorum pro anima est et idcirco non debes animam co-

12:24 Thou shalt not eat it; thou shalt pour it upon the medere cum carnibus 12:24 sed super terram fundes quasi

earth as water.

12:25 Thou shalt not eat it; that it may go well with thee, aquam 12:25 ut sit tibi bene et filiis tuis post te cum feceris

and with thy children after thee, when thou shalt do that which is right in the sight of the LORD.

quod placet in conspectu Domini 12:26 quae autem sancti-

12:26 Only thy holy things which thou hast, and thy vows, thou shalt take, and go unto the place which the LORD shall choose:

ficaueris et uoueris Domino tolles et uenies ad locum quem

12:27 And thou shalt offer thy burnt offerings, the flesh elegerit Dominus 12:27 et offeres oblationes tuas carnem

and the blood, upon the altar of the LORD thy God: and the blood of thy sacrifices shall be poured out upon the

altar of the LORD thy God, and thou shalt eat the flesh. et sanguinem super altare Domini Dei tui sanguinem hos-12:28 Observe and hear all these words which I command tiarum fundes in altari carnibus autem ipse uesceris 12:28

thee, that it may go well with thee, and with thy children after thee for ever, when thou doest that which is good and

right in the sight of the LORD thy God. obserua et audi omnia quae ego praecipio tibi ut bene sit tibi

et filiis tuis post te in sempiternum cum feceris quod bonum

12:29 When the LORD thy God shall cut off the nations est et placitum in conspectu Domini Dei tui 12:29 quando

from before thee, whither thou goest to possess them, and thou succeedest them, and dwellest in their land;

disperderit Dominus Deus tuus ante faciem tuam gentes ad quas ingredieris possidendas et possederis eas atque habita-

12:30 Take heed to thyself that thou be not snared by fol- ueris in terra earum 12:30 caue ne imiteris eas postquam te

lowing them, after that they be destroyed from before thee; and that thou enquire not after their gods, saying, How did

these nations serve their gods? even so will I do likewise. fuerint introeunte subuersae et requiras caerimonias earum

dicens sicut coluerunt gentes istae deos suos ita et ego co-

12:31 Thou shalt not do so unto the LORD thy God: for lam 12:31 non facies similiter Domino Deo tuo omnes enim

every abomination to the LORD, which he hateth, have they done unto their gods; for even their sons and their

daughters they have burnt in the fire to their gods. abominationes quas auersatur Dominus fecerunt diis suis of-12:32 What thing soever I command you, observe to do it: ferentes filios et filias et conburentes igne 12:32 quod prae-

thou shalt not add thereto, nor diminish from it.

cipio tibi hoc tantum facito Domino nec addas quicquam nec minuas

13:1 If there arise among you a prophet, or a dreamer of 13:1 si surrexerit in medio tui prophetes aut qui somnium

dreams, and giveth thee a sign or a wonder,

13:2 And the sign or the wonder come to pass, whereof he uidisse se dicat et praedixerit signum atque portentum 13:2

spake unto thee, saying, Let us go after other gods, which thou hast not known, and let us serve them;

et euenerit quod locutus est et dixerit tibi eamus et sequamur

13:3 Thou shalt not hearken unto the words of that deos alienos quos ignoras et seruiamus eis 13:3 non audies

prophet, or that dreamer of dreams: for the LORD your God proveth you, to know whether ye love the LORD your

God with all your heart and with all your soul. uerba prophetae illius aut somniatoris quia temptat uos Do-

minus Deus uester ut palam fiat utrum diligatis eum an non

13:4 Ye shall walk after the LORD your God, and fear in toto corde et in tota anima uestra 13:4 Dominum Deum

him, and keep his commandments, and obey his voice, and ye shall serve him, and cleave unto him.

uestrum sequimini et ipsum timete mandata illius custodite

13:5 And that prophet, or that dreamer of dreams, shall et audite uocem eius ipsi seruietis et ipsi adherebitis 13:5

be put to death; because he hath spoken to turn you away from the LORD your God, which brought you out of the land of Egypt, and redeemed you out of the house of bondage, to thrust thee out of the way which the LORD thy God commanded thee to walk in. So shalt thou put the evil away from the midst of thee.

propheta autem ille aut fictor somniorum interficietur quia locutus est ut uos auerteret a Domino Deo uestro qui eduxit uos de terra Aegypti et redeeit de domo seruitutis ut errare

Liber Deuteronomii

te faceret de uia quam tibi praecepit Dominus Deus tuus et auferes malum de medio tui 13:6 si tibi uoluerit persuadere 13:6 If thy brother, the son of thy mother, or thy son, or thy

daughter, or the wife of thy bosom, or thy friend, which is as thine own soul, entice thee secretly, saying, Let us go and serve other gods, which thou hast not known, thou, nor thy fathers;

frater tuus filius matris tuae aut filius tuus uel filia siue uxor quae est in sinu tuo aut amicus quem diligis ut animam tuam clam dicens eamus et seruiamus diis alienis quos ignoras tu et patres tui 13:7 cunctarum in circuitu gentium quae iuxta 13:7 Namely, of the gods of the people which are round

about you, nigh unto thee, or far off from thee, from the one end of the earth even unto the other end of the earth;

uel procul sunt ab initio usque ad finem terrae 13:8 non ad-

13:8 Thou shalt not consent unto him, nor hearken unto him; neither shall thine eye pity him, neither shalt thou spare, neither shalt thou conceal him:

quiescas ei nec audias neque parcat ei oculus tuus ut misere-

aris et occultes eum 13:9 sed statim interficies sit primum 13:9 But thou shalt surely kill him; thine hand shall be first upon him to put him to death, and afterwards the hand of all the people.

manus tua super eum et post te omnis populus mittat manum 13:10 lapidibus obrutus necabitur quia uoluit te abstrahere a 13:10 And thou shalt stone him with stones, that he die;

because he hath sought to thrust thee away from the LORD thy God, which brought thee out of the land of Egypt, from

Domino Deo tuo qui eduxit te de terra Aegypti de domo ser- the house of bondage. uitutis 13:11 ut omnis Israhel audiens timeat et nequaquam 13:11 And all Israel shall hear, and fear, and shall do no

more any such wickedness as this is among you.

ultra faciat quippiam huius rei simile 13:12 si audieris in 13:12 If thou shalt hear say in one of thy cities, which the

LORD thy God hath given thee to dwell there, saying,

una urbium tuarum quas Dominus Deus tuus dabit tibi ad habitandum dicentes aliquos 13:13 egressi sunt filii Belial 13:13 Certain men, the children of Belial, are gone out

from among you, and have withdrawn the inhabitants of their city, saying, Let us go and serve other gods, which

de medio tui et auerterunt habitatores urbis tuae atque dixe- ye have not known; runt eamus et seruiamus diis alienis quos ignoratis 13:14 13:14 Then shalt thou enquire, and make search, and ask

diligently; and, behold, if it be truth, and the thing certain, that such abomination is wrought among you;

quaere sollicite et diligenter rei ueritate perspecta si inueneris certum esse quod dicitur et abominationem hanc opere perpetratam 13:15 statim percuties habitatores urbis illius 13:15 Thou shalt surely smite the inhabitants of that city

with the edge of the sword, destroying it utterly, and all that is therein, and the cattle thereof, with the edge of the

in ore gladii et delebis eam omniaque quae in illa sunt usque sword. ad pecora 13:16 quicquid etiam supellectilis fuerit congre- 13:16 And thou shalt gather all the spoil of it into the midst

of the street thereof, and shalt burn with fire the city, and all the spoil thereof every whit, for the LORD thy God:

gabis in medium platearum eius et cum ipsa ciuitate succen- and it shall be an heap for ever; it shall not be built again. des ita ut uniuersa consumas Domino Deo tuo et sit tumulus sempiternus non aedificabitur amplius 13:17 et non adhe- 13:17 And there shall cleave nought of the cursed thing to

thine hand: that the LORD may turn from the fierceness of his anger, and shew thee mercy, and have compassion upon thee, and multiply thee, as he hath sworn unto thy fathers;

rebit de illo anathemate quicquam in manu tua ut auertatur Dominus ab ira furoris sui et misereatur tui multiplicetque te sicut iurauit patribus tuis 13:18 quando audieris uocem 13:18 When thou shalt hearken to the voice of the LORD

thy God, to keep all his commandments which I command thee this day, to do that which is right in the eyes of the

Domini Dei tui custodiens omnia praecepta eius quae ego LORD thy God. praecipio tibi hodie ut facias quod placitum est in conspectu Domini Dei tui

14:1 filii estote Domini Dei uestri non uos incidetis nec fa- 14:1 Ye are the children of the LORD your God: ye shall

not cut yourselves, nor make any baldness between your eyes for the dead.

Biblia Sacra Vulgata

14:2 For thou art an holy people unto the LORD thy God, cietis caluitium super mortuo 14:2 quoniam populus sanc-

and the LORD hath chosen thee to be a peculiar people unto himself, above all the nations that are upon the earth.

tus es Domino Deo tuo et te elegit ut sis ei in populum pecu-

14:3 Thou shalt not eat any abominable thing. liarem de cunctis gentibus quae sunt super terram 14:3 ne 14:4 These are the beasts which ye shall eat: the ox, the comedatis quae inmunda sunt 14:4 hoc est animal quod co-

sheep, and the goat, 14:5 The hart, and the roebuck, and the fallow deer, and medere debetis bouem et ouem et capram 14:5 ceruum capthe wild goat, and the pygarg, and the wild ox, and the chamois.

ream bubalum tragelaphum pygargon orygem cameloparda-

14:6 And every beast that parteth the hoof, and cleaveth lum 14:6 omne animal quod in duas partes ungulam findit the cleft into two claws, and cheweth the cud among the beasts, that ye shall eat.

et ruminat comedetis 14:7 de his autem quae ruminant et

14:7 Nevertheless these ye shall not eat of them that chew the cud, or of them that divide the cloven hoof; as the camel, and the hare, and the coney: for they chew the cud, but divide not the hoof; therefore they are unclean unto you.

ungulam non findunt haec comedere non debetis camelum leporem choerogyllium quia ruminant et non diuidunt un-

14:8 And the swine, because it divideth the hoof, yet gulam inmunda erunt uobis 14:8 sus quoque quoniam di-

cheweth not the cud, it is unclean unto you: ye shall not eat of their flesh, nor touch their dead carcase.

uidit ungulam et non ruminat inmunda erit carnibus eorum

14:9 These ye shall eat of all that are in the waters: all that non uescemini et cadauera non tangetis 14:9 haec comehave fins and scales shall ye eat:

detis ex omnibus quae morantur in aquis quae habent pin-

14:10 And whatsoever hath not fins and scales ye may not nulas et squamas comedite 14:10 quae absque pinnulis et

eat; it is unclean unto you.

14:11 Of all clean birds ye shall eat. squamis sunt ne comedatis quia inmunda sunt 14:11 omnes 14:12 But these are they of which ye shall not eat: the aues mundas comedite 14:12 inmundas ne comedatis aqui-

eagle, and the ossifrage, and the ospray, 14:13 And the glede, and the kite, and the vulture after his lam scilicet et grypem et alietum 14:13 ixon et uulturem kind,

14:14 And every raven after his kind, ac miluum iuxta genus suum 14:14 et omne coruini gene-14:15 And the owl, and the night hawk, and the cuckow, ris 14:15 strutionem ac noctuam et larum atque accipitrem

and the hawk after his kind,

14:16 The little owl, and the great owl, and the swan, iuxta genus suum 14:16 herodium et cycnum et ibin 14:17

14:17 And the pelican, and the gier eagle, and the cormorant,

ac mergulum porphirionem et nycticoracem 14:18 onocro-

14:18 And the stork, and the heron after her kind, and the lapwing, and the bat.

talum et charadrium singula in genere suo upupam quoque et

14:19 And every creeping thing that flieth is unclean unto uespertilionem 14:19 et omne quod reptat et pinnulas habet

you: they shall not be eaten.

14:20 But of all clean fowls ye may eat. inmundum erit nec comedetur 14:20 omne quod mundum

14:21 Ye shall not eat of anything that dieth of itself: thou est comedite 14:21 quicquid morticinum est ne uescamini shalt give it unto the stranger that is in thy gates, that he may eat it; or thou mayest sell it unto an alien: for thou art an holy people unto the LORD thy God. Thou shalt not seethe a kid in his mother’s milk.

ex eo peregrino qui intra portas tuas est da ut comedat aut uende ei quia tu populus sanctus Domini Dei tui es non co-

14:22 Thou shalt truly tithe all the increase of thy seed, ques hedum in lacte matris suae 14:22 decimam partem sethat the field bringeth forth year by year.

parabis de cunctis frugibus tuis quae nascuntur in terra per

14:23 And thou shalt eat before the LORD thy God, in annos singulos 14:23 et comedes in conspectu Domini Dei

the place which he shall choose to place his name there, the tithe of thy corn, of thy wine, and of thine oil, and the firstlings of thy herds and of thy flocks; that thou mayest learn to fear the LORD thy God always.

tui in loco quem elegerit ut in eo nomen illius inuocetur decimam frumenti tui et uini et olei et primogenita de armentis

Liber Deuteronomii

et ouibus tuis ut discas timere Dominum Deum tuum omni tempore 14:24 cum autem longior fuerit uia et locus quem 14:24 And if the way be too long for thee, so that thou art

not able to carry it; or if the place be too far from thee, which the LORD thy God shall choose to set his name

elegerit Dominus Deus tuus tibique benedixerit nec potueris there, when the LORD thy God hath blessed thee: ad eum haec cuncta portare 14:25 uendes omnia et in pre- 14:25 Then shalt thou turn it into money, and bind up the

money in thine hand, and shalt go unto the place which the LORD thy God shall choose:

tium rediies portabisque manu tua et proficisceris ad locum quem elegerit Dominus Deus tuus 14:26 et emes ex eadem 14:26 And thou shalt bestow that money for whatsoever

thy soul lusteth after, for oxen, or for sheep, or for wine, or for strong drink, or for whatsoever thy soul desireth: and thou shalt eat there before the LORD thy God, and thou shalt rejoice, thou, and thine household,

pecunia quicquid tibi placuerit siue ex armentis siue ex ouibus uinum quoque et siceram et omne quod desiderat anima tua et comedes coram Domino Deo tuo et epulaberis tu et domus tua 14:27 et Leuita qui intra portas tuas est caue ne 14:27 And the Levite that is within thy gates; thou shalt

not forsake him; for he hath no part nor inheritance with thee.

derelinquas eum quia non habet aliam partem in possessione tua 14:28 anno tertio separabis aliam decimam ex omnibus 14:28 At the end of three years thou shalt bring forth all

the tithe of thine increase the same year, and shalt lay it up within thy gates:

quae nascuntur tibi eo tempore et repones intra ianuas tuas 14:29 uenietque Leuites qui aliam non habet partem nec pos- 14:29 And the Levite, (because he hath no part nor inher-

itance with thee,) and the stranger, and the fatherless, and the widow, which are within thy gates, shall come, and shall eat and be satisfied; that the LORD thy God may bless thee in all the work of thine hand which thou doest.

sessionem tecum et peregrinus et pupillus ac uidua qui intra portas tuas sunt et comedent et saturabuntur ut benedicat tibi Dominus Deus tuus in cunctis operibus manuum tuarum quae feceris

15:1 septimo anno facies remissionem 15:2 quae hoc or- 15:1 At the end of every seven years thou shalt make a

release.

15:2 And this is the manner of the release: Every creditor that lendeth ought unto his neighbour shall release it; he shall not exact it of his neighbour, or of his brother; because it is called the LORD’s release.

dine celebrabitur cui debetur aliquid ab amico uel proximo ac fratre suo repetere non poterit quia annus remissionis est Domini 15:3 a peregrino et aduena exiges ciuem et propin- 15:3 Of a foreigner thou mayest exact it again: but that

which is thine with thy brother thine hand shall release;

quum repetendi non habes potestatem 15:4 et omnino indi- 15:4 Save when there shall be no poor among you; for the

LORD shall greatly bless thee in the land which the LORD thy God giveth thee for an inheritance to possess it:

gens et mendicus non erit inter uos ut benedicat tibi Dominus in terra quam traditurus est tibi in possessionem 15:5 15:5 Only if thou carefully hearken unto the voice of the

LORD thy God, to observe to do all these commandments which I command thee this day.

si tamen audieris uocem Domini Dei tui et custodieris uniuersa quae iussit et quae ego hodie praecipio tibi benedicet tibi ut pollicitus est 15:6 fenerabis gentibus multis et 15:6 For the LORD thy God blesseth thee, as he promised

thee: and thou shalt lend unto many nations, but thou shalt not borrow; and thou shalt reign over many nations, but

ipse a nullo accipies mutuum dominaberis nationibus pluri- they shall not reign over thee. mis et tui nemo dominabitur 15:7 si unus de fratribus tuis 15:7 If there be among you a poor man of one of thy

brethren within any of thy gates in thy land which the LORD thy God giveth thee, thou shalt not harden thine

qui morantur intra portas ciuitatis tuae in terra quam Domi- heart, nor shut thine hand from thy poor brother: nus Deus tuus daturus est tibi ad paupertatem uenerit non obdurabis cor tuum nec contrahes manum 15:8 sed aperies 15:8 But thou shalt open thine hand wide unto him, and

shalt surely lend him sufficient for his need, in that which he wanteth.

Biblia Sacra Vulgata

eam pauperi et dabis mutuum quod eum indigere perspexe-

15:9 Beware that there be not a thought in thy wicked ris 15:9 caue ne forte subripiat tibi impia cogitatio et di-

heart, saying, The seventh year, the year of release, is at hand; and thine eye be evil against thy poor brother, and thou givest him nought; and he cry unto the LORD against thee, and it be sin unto thee.

cas in corde tuo adpropinquat septimus annus remissionis et auertas oculos a paupere fratre tuo nolens ei quod postulat mutuum commodare ne clamet contra te ad Dominum et fiat

15:10 Thou shalt surely give him, and thine heart shall not tibi in peccatum 15:10 sed dabis ei nec ages quippiam calbe grieved when thou givest unto him: because that for this thing the LORD thy God shall bless thee in all thy works,

and in all that thou puttest thine hand unto. lide in eius necessitatibus subleuandis ut benedicat tibi Do-

minus Deus tuus in omni tempore et in cunctis ad quae ma-

15:11 For the poor shall never cease out of the land: there- num miseris 15:11 non deerunt pauperes in terra habitatio-

fore I command thee, saying, Thou shalt open thine hand wide unto thy brother, to thy poor, and to thy needy, in thy

land. nis tuae idcirco ego praecipio tibi ut aperias manum fratri tuo 15:12 And if thy brother, an Hebrew man, or an Hebrew egeno et pauperi qui tecum uersatur in terra 15:12 cum tibi

woman, be sold unto thee, and serve thee six years; then in the seventh year thou shalt let him go free from thee.

uenditus fuerit frater tuus hebraeus aut hebraea et sex annis

15:13 And when thou sendest him out free from thee, thou seruierit tibi in septimo anno dimittes eum liberum 15:13 et

shalt not let him go away empty:

quem libertate donaueris nequaquam uacuum abire patieris

15:14 Thou shalt furnish him liberally out of thy flock, and 15:14 sed dabis uiaticum de gregibus et de area et torculari

out of thy floor, and out of thy winepress: of that wherewith the LORD thy God hath blessed thee thou shalt give

unto him. tuo quibus Dominus Deus tuus benedixerit tibi 15:15 me-

15:15 And thou shalt remember that thou wast a bondman in the land of Egypt, and the LORD thy God redeemed thee: therefore I command thee this thing to day.

mento quod et ipse seruieris in terra Aegypti et liberauerit te Dominus Deus tuus et idcirco ego nunc praecipiam tibi

15:16 And it shall be, if he say unto thee, I will not go 15:16 sin autem dixerit nolo egredi eo quod diligat te et do-

away from thee; because he loveth thee and thine house, because he is well with thee;

mum tuam et bene sibi apud te esse sentiat 15:17 adsumes

15:17 Then thou shalt take an aul, and thrust it through his ear unto the door, and he shall be thy servant for ever. And also unto thy maidservant thou shalt do likewise.

subulam et perforabis aurem eius in ianua domus tuae et seruiet tibi usque in aeternum ancillae quoque similiter facies

15:18 It shall not seem hard unto thee, when thou sendest 15:18 non auertes ab eis oculos tuos quando dimiseris eos li-

him away free from thee; for he hath been worth a double hired servant to thee, in serving thee six years: and the

LORD thy God shall bless thee in all that thou doest. beros quoniam iuxta mercedem mercennarii per sex annos

seruiuit tibi ut benedicat tibi Dominus Deus tuus in cunctis

15:19 All the firstling males that come of thy herd and of operibus quae agis 15:19 de primogenitis quae nascuntur in

thy flock thou shalt sanctify unto the LORD thy God: thou shalt do no work with the firstling of thy bullock, nor shear

the firstling of thy sheep. armentis et ouibus tuis quicquid sexus est masculini sanctifi-

cabis Domino Deo tuo non operaberis in primogenito bouis

15:20 Thou shalt eat it before the LORD thy God year by et non tondebis primogenita ouium 15:20 in conspectu Do-

year in the place which the LORD shall choose, thou and thy household.

mini Dei tui comedes ea per annos singulos in loco quem

15:21 And if there be any blemish therein, as if it be lame, elegerit Dominus tu et domus tua 15:21 sin autem habueor blind, or have any ill blemish, thou shalt not sacrifice it unto the LORD thy God.

rit maculam et uel claudum fuerit uel caecum aut in aliqua parte deforme uel debile non immolabitur Domino Deo tuo

Liber Deuteronomii

15:22 sed intra portas urbis tuae comedes illud tam mundus 15:22 Thou shalt eat it within thy gates: the unclean and

the clean person shall eat it alike, as the roebuck, and as the hart.

quam inmundus similiter uescentur eis quasi caprea et ceruo

15:23 hoc solum obseruabis ut sanguinem eorum non come- 15:23 Only thou shalt not eat the blood thereof; thou shalt pour it upon the ground as water.

das sed effundas in terram quasi aquam

16:1 obserua mensem nouarum frugum et uerni primum 16:1 Observe the month of Abib, and keep the passover unto the LORD thy God: for in the month of Abib the LORD thy God brought thee forth out of Egypt by night.

temporis ut facias phase Domino Deo tuo quoniam in isto

mense eduxit te Dominus Deus tuus de Aegypto nocte 16:2 16:2 Thou shalt therefore sacrifice the passover unto the LORD thy God, of the flock and the herd, in the place which the LORD shall choose to place his name there.

immolabisque phase Domino Deo tuo de ouibus et de bubus in loco quem elegerit Dominus Deus tuus ut habitet no-

men eius ibi 16:3 non comedes in eo panem fermentatum 16:3 Thou shalt eat no leavened bread with it; seven days shalt thou eat unleavened bread therewith, even the bread of affliction; for thou camest forth out of the land of Egypt in haste: that thou mayest remember the day when thou camest forth out of the land of Egypt all the days of thy life.

septem diebus comedes absque fermento adflictionis panem quoniam in pauore egressus es de Aegypto ut memineris diei egressionis tuae de Aegypto omnibus diebus uitae tuae 16:4 16:4 And there shall be no leavened bread seen with thee

in all thy coast seven days; neither shall there any thing of the flesh, which thou sacrificedst the first day at even,

non apparebit fermentum in omnibus terminis tuis septem remain all night until the morning. diebus et non manebit de carnibus eius quod immolatum est uesperi in die primo mane 16:5 non poteris immolare phase 16:5 Thou mayest not sacrifice the passover within any of

thy gates, which the LORD thy God giveth thee:

in qualibet urbium tuarum quas Dominus Deus tuus daturus est tibi 16:6 sed in loco quem elegerit Dominus Deus tuus 16:6 But at the place which the LORD thy God shall

choose to place his name in, there thou shalt sacrifice the passover at even, at the going down of the sun, at the sea-

ut habitet nomen eius ibi immolabis phase uesperi ad solis son that thou camest forth out of Egypt.

occasum quando egressus es de Aegypto 16:7 et coques et 16:7 And thou shalt roast and eat it in the place which the LORD thy God shall choose: and thou shalt turn in the morning, and go unto thy tents.

comedes in loco quem elegerit Dominus Deus tuus maneque consurgens uades in tabernacula tua 16:8 sex diebus 16:8 Six days thou shalt eat unleavened bread: and on the

seventh day shall be a solemn assembly to the LORD thy God: thou shalt do no work therein.

comedes azyma et in die septimo quia collecta est Domini Dei tui non facies opus 16:9 septem ebdomadas numera- 16:9 Seven weeks shalt thou number unto thee: begin to

number the seven weeks from such time as thou beginnest to put the sickle to the corn.

bis tibi ab ea die qua falcem in segetem miseris 16:10 et

16:10 And thou shalt keep the feast of weeks unto the LORD thy God with a tribute of a freewill offering of thine hand, which thou shalt give unto the LORD thy God, according as the LORD thy God hath blessed thee:

celebrabis diem festum ebdomadarum Domino Deo tuo oblationem spontaneam manus tuae quam offeres iuxta benedictionem Domini Dei tui 16:11 et epulaberis coram Do- 16:11 And thou shalt rejoice before the LORD thy God,

thou, and thy son, and thy daughter, and thy manservant, and thy maidservant, and the Levite that is within thy gates, and the stranger, and the fatherless, and the widow, that are among you, in the place which the LORD thy God hath chosen to place his name there.

mino Deo tuo tu et filius tuus et filia tua et seruus tuus et ancilla et Leuites qui est intra portas tuas et aduena ac pupillus et uidua qui morantur uobiscum in loco quem elegerit Dominus Deus tuus ut habitet nomen eius ibi 16:12 et re- 16:12 And thou shalt remember that thou wast a bondman

in Egypt: and thou shalt observe and do these statutes.

cordaberis quoniam seruus fueris in Aegypto custodiesque

Biblia Sacra Vulgata

16:13 Thou shalt observe the feast of tabernacles seven ac facies quae praecepta sunt 16:13 sollemnitatem quoque

days, after that thou hast gathered in thy corn and thy wine:

tabernaculorum celebrabis per septem dies quando college-

16:14 And thou shalt rejoice in thy feast, thou, and thy ris de area et torculari fruges tuas 16:14 et epulaberis in fe-

son, and thy daughter, and thy manservant, and thy maidservant, and the Levite, the stranger, and the fatherless,

and the widow, that are within thy gates. stiuitate tua tu et filius tuus et filia et seruus tuus et ancilla

Leuites quoque et aduena et pupillus ac uidua qui intra por-

16:15 Seven days shalt thou keep a solemn feast unto the tas tuas sunt 16:15 septem diebus Domino Deo tuo festa LORD thy God in the place which the LORD shall choose: because the LORD thy God shall bless thee in all thine increase, and in all the works of thine hands, therefore thou shalt surely rejoice.

celebrabis in loco quem elegerit Dominus benedicetque tibi Dominus Deus tuus in cunctis frugibus tuis et in omni opere

16:16 Three times in a year shall all thy males appear be- manuum tuarum erisque in laetitia 16:16 tribus uicibus per

fore the LORD thy God in the place which he shall choose; in the feast of unleavened bread, and in the feast of weeks, and in the feast of tabernacles: and they shall not appear before the LORD empty:

annum apparebit omne masculinum tuum in conspectu Domini Dei tui in loco quem elegerit in sollemnitate azymorum et in sollemnitate ebdomadarum et in sollemnitate taberna-

16:17 Every man shall give as he is able, according to the culorum non apparebit ante Dominum uacuus 16:17 sed ofblessing of the LORD thy God which he hath given thee.

feret unusquisque secundum quod habuerit iuxta benedictio-

16:18 Judges and officers shalt thou make thee in all thy nem Domini Dei sui quam dederit ei 16:18 iudices et magates, which the LORD thy God giveth thee, throughout thy tribes: and they shall judge the people with just judg-

ment. gistros constitues in omnibus portis tuis quas Dominus Deus

tuus dederit tibi per singulas tribus tuas ut iudicent populum

16:19 Thou shalt not wrest judgment; thou shalt not re- iusto iudicio 16:19 nec in alteram partem declinent non ac-

spect persons, neither take a gift: for a gift doth blind the eyes of the wise, and pervert the words of the righteous.

cipies personam nec munera quia munera excaecant oculos

16:20 That which is altogether just shalt thou follow, that sapientium et mutant uerba iustorum 16:20 iuste quod iusthou mayest live, and inherit the land which the LORD thy God giveth thee.

tum est persequeris ut uiuas et possideas terram quam Do-

16:21 Thou shalt not plant thee a grove of any trees near minus Deus tuus dederit tibi 16:21 non plantabis lucum et

unto the altar of the LORD thy God, which thou shalt make thee.

omnem arborem iuxta altare Domini Dei tui 16:22 nec fa-

16:22 Neither shalt thou set thee up any image; which the LORD thy God hateth.

cies tibi atque constitues statuam quae odit Dominus Deus tuus

17:1 Thou shalt not sacrifice unto the LORD thy God any 17:1 non immolabis Domino Deo tuo bouem et ouem in bullock, or sheep, wherein is blemish, or any evilfavouredness: for that is an abomination unto the LORD thy God.

quo est macula aut quippiam uitii quia abominatio est Do-

17:2 If there be found among you, within any of thy gates mini Dei tui 17:2 cum repperti fuerint apud te intra unam

which the LORD thy God giveth thee, man or woman, that hath wrought wickedness in the sight of the LORD

thy God, in transgressing his covenant, portarum tuarum quas Dominus Deus tuus dabit tibi uir aut

mulier qui faciant malum in conspectu Domini Dei tui et

17:3 And hath gone and served other gods, and wor- transgrediantur pactum illius 17:3 ut uadant et seruiant diis

shipped them, either the sun, or moon, or any of the host of heaven, which I have not commanded;

alienis et adorent eos solem et lunam et omnem militiam

17:4 And it be told thee, and thou hast heard of it, and caeli quae non praecepi 17:4 et hoc tibi fuerit nuntiatum

enquired diligently, and, behold, it be true, and the thing certain, that such abomination is wrought in Israel:

Liber Deuteronomii

audiensque inquisieris diligenter et uerum esse reppereris et abominatio facta est in Israhel 17:5 educes uirum ac mulie- 17:5 Then shalt thou bring forth that man or that woman,

which have committed that wicked thing, unto thy gates, even that man or that woman, and shalt stone them with

rem qui rem sceleratissimam perpetrarunt ad portas ciuitatis stones, till they die. tuae et lapidibus obruentur 17:6 in ore duorum aut trium te- 17:6 At the mouth of two witnesses, or three witnesses,

shall he that is worthy of death be put to death; but at the mouth of one witness he shall not be put to death.

stium peribit qui interficietur nemo occidatur uno contra se dicente testimonium 17:7 manus testium prima interficiet 17:7 The hands of the witnesses shall be first upon him to

put him to death, and afterward the hands of all the people. So thou shalt put the evil away from among you.

eum et manus reliqui populi extrema mittetur ut auferas ma-

lum de medio tui 17:8 si difficile et ambiguum apud te iu- 17:8 If there arise a matter too hard for thee in judgment, between blood and blood, between plea and plea, and between stroke and stroke, being matters of controversy within thy gates: then shalt thou arise, and get thee up into the place which the LORD thy God shall choose;

dicium esse perspexeris inter sanguinem et sanguinem causam et causam lepram et non lepram et iudicum intra portas tuas uideris uerba uariari surge et ascende ad locum quem elegerit Dominus Deus tuus 17:9 ueniesque ad sacerdotes 17:9 And thou shalt come unto the priests the Levites, and

unto the judge that shall be in those days, and enquire; and they shall shew thee the sentence of judgment:

leuitici generis et ad iudicem qui fuerit illo tempore quaeresque ab eis qui indicabunt tibi iudicii ueritatem 17:10 et 17:10 And thou shalt do according to the sentence, which

they of that place which the LORD shall choose shall shew thee; and thou shalt observe to do according to all that they

facies quodcumque dixerint qui praesunt loco quem elege- inform thee: rit Dominus et docuerint te 17:11 iuxta legem eius sequeris 17:11 According to the sentence of the law which they

shall teach thee, and according to the judgment which they shall tell thee, thou shalt do: thou shalt not decline from the sentence which they shall shew thee, to the right hand, nor to the left.

sententiam eorum nec declinabis ad dextram uel ad sinistram 17:12 qui autem superbierit nolens oboedire sacerdotis impe- 17:12 And the man that will do presumptuously, and will

not hearken unto the priest that standeth to minister there before the LORD thy God, or unto the judge, even that

rio qui eo tempore ministrat Domino Deo tuo et decreto iu- man shall die: and thou shalt put away the evil from Israel.

dicis morietur homo ille et auferes malum de Israhel 17:13 17:13 And all the people shall hear, and fear, and do no

more presumptuously.

cunctusque populus audiens timebit ut nullus deinceps intumescat superbia 17:14 cum ingressus fueris terram quam 17:14 When thou art come unto the land which the LORD

thy God giveth thee, and shalt possess it, and shalt dwell therein, and shalt say, I will set a king over me, like as all

Dominus Deus tuus dabit tibi et possederis eam habitaueris- the nations that are about me; que in illa et dixeris constituam super me regem sicut habent omnes per circuitum nationes 17:15 eum constitues quem 17:15 Thou shalt in any wise set him king over thee, whom

the LORD thy God shall choose: one from among thy brethren shalt thou set king over thee: thou mayest not

Dominus Deus tuus elegerit de numero fratrum tuorum non set a stranger over thee, which is not thy brother. poteris alterius gentis hominem regem facere qui non sit frater tuus 17:16 cumque fuerit constitutus non multiplicabit 17:16 But he shall not multiply horses to himself, nor

cause the people to return to Egypt, to the end that he should multiply horses: forasmuch as the LORD hath said

sibi equos nec reducet populum in Aegyptum equitatus nu- unto you, Ye shall henceforth return no more that way. mero subleuatus praesertim cum Dominus praeceperit uobis

ut nequaquam amplius per eandem uiam reuertamini 17:17 17:17 Neither shall he multiply wives to himself, that his heart turn not away: neither shall he greatly multiply to himself silver and gold.

non habebit uxores plurimas quae inliciant animum eius neque argenti et auri inmensa pondera 17:18 postquam autem 17:18 And it shall be, when he sitteth upon the throne of

his kingdom, that he shall write him a copy of this law in a book out of that which is before the priests the Levites:

Biblia Sacra Vulgata

sederit in solio regni sui describet sibi deuteronomium legis huius in uolumine accipiens exemplar a sacerdotibus leui-

17:19 And it shall be with him, and he shall read therein all ticae tribus 17:19 et habebit secum legetque illud omnibus

the days of his life: that he may learn to fear the LORD his God, to keep all the words of this law and these statutes,

to do them: diebus uitae suae ut discat timere Dominum Deum suum et

custodire uerba et caerimonias eius quae lege praecepta sunt

17:20 That his heart be not lifted up above his brethren, 17:20 nec eleuetur cor eius in superbiam super fratres suos

and that he turn not aside from the commandment, to the right hand, or to the left: to the end that he may prolong his days in his kingdom, he, and his children, in the midst of Israel.

neque declinet in partem dextram uel sinistram ut longo tempore regnet ipse et filii eius super Israhel

18:1 The priests the Levites, and all the tribe of Levi, shall 18:1 non habebunt sacerdotes et Leuitae et omnes qui de have no part nor inheritance with Israel: they shall eat the offerings of the LORD made by fire, and his inheritance.

eadem tribu sunt partem et hereditatem cum reliquo Israhel

18:2 Therefore shall they have no inheritance among their quia sacrificia Domini et oblationes eius comedent 18:2 et brethren: the LORD is their inheritance, as he hath said unto them.

nihil aliud accipient de possessione fratrum suorum Domi-

18:3 And this shall be the priest’s due from the people, nus enim ipse est hereditas eorum sicut locutus est illis 18:3

from them that offer a sacrifice, whether it be ox or sheep; and they shall give unto the priest the shoulder, and the

two cheeks, and the maw. hoc erit iudicium sacerdotum a populo et ab his qui offerunt

uictimas siue bouem siue ouem immolauerint dabunt sacer-

18:4 The firstfruit also of thy corn, of thy wine, and of doti armum ac uentriculum 18:4 primitias frumenti uini et

thine oil, and the first of the fleece of thy sheep, shalt thou give him.

olei et lanarum partem ex ouium tonsione 18:5 ipsum enim

18:5 For the LORD thy God hath chosen him out of all thy tribes, to stand to minister in the name of the LORD, him and his sons for ever.

elegit Dominus Deus tuus de cunctis tribubus tuis ut stet et ministret nomini Domini ipse et filii eius in sempiternum

18:6 And if a Levite come from any of thy gates out of all 18:6 si exierit Leuites de una urbium tuarum ex omni Isra-

Israel, where he sojourned, and come with all the desire of his mind unto the place which the LORD shall choose;

hel in qua habitat et uoluerit uenire desiderans locum quem

18:7 Then he shall minister in the name of the LORD his elegerit Dominus 18:7 ministrabit in nomine Dei sui sic-

God, as all his brethren the Levites do, which stand there before the LORD.

ut omnes fratres eius Leuitae qui stabunt eo tempore coram

18:8 They shall have like portions to eat, beside that which Domino 18:8 partem ciborum eandem accipiet quam et cecometh of the sale of his patrimony.

teri excepto eo quod in urbe sua ex paterna ei successione

18:9 When thou art come into the land which the LORD debetur 18:9 quando ingressus fueris terram quam Domi-

thy God giveth thee, thou shalt not learn to do after the abominations of those nations.

nus Deus tuus dabit tibi caue ne imitari uelis abominationes

18:10 There shall not be found among you any one that illarum gentium 18:10 nec inueniatur in te qui lustret filium

maketh his son or his daughter to pass through the fire, or that useth divination, or an observer of times, or an en-

chanter, or a witch. suum aut filiam ducens per ignem aut qui ariolos sciscite-

18:11 Or a charmer, or a consulter with familiar spirits, or tur et obseruet somnia atque auguria ne sit maleficus 18:11 a wizard, or a necromancer.

ne incantator ne pythones consulat ne diuinos et quaerat a

18:12 For all that do these things are an abomination unto mortuis ueritatem 18:12 omnia enim haec abominatur Dothe LORD: and because of these abominations the LORD thy God doth drive them out from before thee.

minus et propter istiusmodi scelera delebit eos in introitu tuo

Liber Deuteronomii

18:13 perfectus eris et absque macula cum Domino Deo tuo 18:13 Thou shalt be perfect with the LORD thy God. 18:14 gentes istae quarum possidebis terram augures et diui- 18:14 For these nations, which thou shalt possess, hear-

kened unto observers of times, and unto diviners: but as for thee, the LORD thy God hath not suffered thee so to

nos audiunt tu autem a Domino Deo tuo aliter institutus es do. 18:15 prophetam de gente tua et de fratribus tuis sicut me su- 18:15 The LORD thy God will raise up unto thee a Prophet

from the midst of thee, of thy brethren, like unto me; unto him ye shall hearken;

scitabit tibi Dominus Deus tuus ipsum audies 18:16 ut pe-

18:16 According to all that thou desiredst of the LORD thy God in Horeb in the day of the assembly, saying, Let me not hear again the voice of the LORD my God, neither let me see this great fire any more, that I die not.

tisti a Domino Deo tuo in Horeb quando contio congregata est atque dixisti ultra non audiam uocem Domini Dei mei et ignem hunc maximum amplius non uidebo ne moriar 18:17 18:17 And the LORD said unto me, They have well spo-

ken that which they have spoken.

et ait Dominus mihi bene omnia sunt locuti 18:18 prophe- 18:18 I will raise them up a Prophet from among their

brethren, like unto thee, and will put my words in his mouth; and he shall speak unto them all that I shall com-

tam suscitabo eis de medio fratrum suorum similem tui et mand him. ponam uerba mea in ore eius loqueturque ad eos omnia quae praecepero illi 18:19 qui autem uerba eius quae loquetur in 18:19 And it shall come to pass, that whosoever will not

hearken unto my words which he shall speak in my name, I will require it of him.

nomine meo audire noluerit ego ultor existam 18:20 pro-

18:20 But the prophet, which shall presume to speak a word in my name, which I have not commanded him to speak, or that shall speak in the name of other gods, even that prophet shall die.

pheta autem qui arrogantia deprauatus uoluerit loqui in nomine meo quae ego non praecepi illi ut diceret aut ex nomine alienorum deorum interficietur 18:21 quod si tacita cogita- 18:21 And if thou say in thine heart, How shall we know

the word which the LORD hath not spoken?

tione responderis quomodo possum intellegere uerbum quod non est locutus Dominus 18:22 hoc habebis signum quod in 18:22 When a prophet speaketh in the name of the LORD,

if the thing follow not, nor come to pass, that is the thing which the LORD hath not spoken, but the prophet hath

nomine Domini propheta ille praedixerit et non euenerit hoc spoken it presumptuously: thou shalt not be afraid of him. Dominus non locutus est sed per tumorem animi sui propheta confinxit et idcirco non timebis eum

19:1 cum disperderit Dominus Deus tuus gentes quarum 19:1 When the LORD thy God hath cut off the nations,

whose land the LORD thy God giveth thee, and thou succeedest them, and dwellest in their cities, and in their

tibi traditurus est terram et possederis eam habitauerisque in houses; urbibus eius et in aedibus 19:2 tres ciuitates separabis tibi 19:2 Thou shalt separate three cities for thee in the midst

of thy land, which the LORD thy God giveth thee to possess it.

in medio terrae quam Dominus Deus tuus dabit tibi in possessionem 19:3 sternens diligenter uiam et in tres aequali- 19:3 Thou shalt prepare thee a way, and divide the coasts

of thy land, which the LORD thy God giveth thee to inherit, into three parts, that every slayer may flee thither.

ter partes totam terrae tuae prouinciam diuides ut habeat e uicino qui propter homicidium profugus est quo possit euadere 19:4 haec erit lex homicidae fugientis cuius uita ser- 19:4 And this is the case of the slayer, which shall flee

thither, that he may live: Whoso killeth his neighbour ignorantly, whom he hated not in time past;

uanda est qui percusserit proximum suum nesciens et qui heri et nudius tertius nullum contra eum habuisse odium conprobatur 19:5 sed abisse simpliciter cum eo in siluam 19:5 As when a man goeth into the wood with his neigh-

bour to hew wood, and his hand fetcheth a stroke with the axe to cut down the tree, and the head slippeth from the helve, and lighteth upon his neighbour, that he die; he shall flee unto one of those cities, and live:

ad ligna caedenda et in succisione lignorum securis fugerit

Biblia Sacra Vulgata

manu ferrumque lapsum de manubrio amicum eius percusserit et occiderit hic ad unam supradictarum urbium confu-

19:6 Lest the avenger of the blood pursue the slayer, while giet et uiuet 19:6 ne forsitan proximus eius cuius effusus est

his heart is hot, and overtake him, because the way is long, and slay him; whereas he was not worthy of death, inas-

much as he hated him not in time past. sanguis dolore stimulatus persequatur et adprehendat eum si longior uia fuerit et percutiat animam eius qui non est reus mortis quia nullum contra eum qui occisus est odium prius

19:7 Wherefore I command thee, saying, Thou shalt sepa- habuisse monstratur 19:7 idcirco praecipio tibi ut tres ciui-

rate three cities for thee.

19:8 And if the LORD thy God enlarge thy coast, as he tates aequalis inter se spatii diuidas 19:8 cum autem dilata-

hath sworn unto thy fathers, and give thee all the land which he promised to give unto thy fathers;

uerit Dominus Deus tuus terminos tuos sicut iurauit patribus tuis et dederit tibi cunctam terram quam eis pollicitus est

19:9 If thou shalt keep all these commandments to do 19:9 si tamen custodieris mandata eius et feceris quae hodie

them, which I command thee this day, to love the LORD thy God, and to walk ever in his ways; then shalt thou add

three cities more for thee, beside these three: praecipio tibi ut diligas Dominum Deum tuum et ambules

in uiis eius omni tempore addes tibi tres alias ciuitates et

19:10 That innocent blood be not shed in thy land, which supradictarum trium urbium numerum duplicabis 19:10 ut

the LORD thy God giveth thee for an inheritance, and so blood be upon thee.

non effundatur sanguis innoxius in medio terrae quam Dominus Deus tuus dabit tibi possidendam nec sis sanguinis

19:11 But if any man hate his neighbour, and lie in wait reus 19:11 si quis autem odio habens proximum suum insi-

for him, and rise up against him, and smite him mortally that he die, and fleeth into one of these cities:

diatus fuerit uitae eius surgensque percusserit illum et mor-

19:12 Then the elders of his city shall send and fetch him tuus fuerit fugeritque ad unam de supradictis urbibus 19:12

thence, and deliver him into the hand of the avenger of blood, that he may die.

mittent seniores ciuitatis illius et arripient eum de loco effugii tradentque in manu proximi cuius sanguis effusus est

19:13 Thine eye shall not pity him, but thou shalt put away et morietur 19:13 nec misereberis eius et auferes innoxium

the guilt of innocent blood from Israel, that it may go well with thee.

sanguinem de Israhel ut bene sit tibi 19:14 non adsumes

19:14 Thou shalt not remove thy neighbour’s landmark, which they of old time have set in thine inheritance, which thou shalt inherit in the land that the LORD thy God giveth thee to possess it.

et transferes terminos proximi tui quos fixerunt priores in possessione tua quam Dominus Deus tuus dabit tibi in terra

19:15 One witness shall not rise up against a man for any quam acceperis possidendam 19:15 non stabit testis unus

iniquity, or for any sin, in any sin that he sinneth: at the mouth of two witnesses, or at the mouth of three witnesses,

shall the matter be established. contra aliquem quicquid illud peccati et facinoris fuerit sed

19:16 If a false witness rise up against any man to testify in ore duorum aut trium testium stabit omne uerbum 19:16 against him that which is wrong;

si steterit testis mendax contra hominem accusans eum prae-

19:17 Then both the men, between whom the controversy uaricationis 19:17 stabunt ambo quorum causa est ante Do-

is, shall stand before the LORD, before the priests and the judges, which shall be in those days;

minum in conspectu sacerdotum et iudicum qui fuerint in

19:18 And the judges shall make diligent inquisition: and, diebus illis 19:18 cumque diligentissime perscrutantes in-

behold, if the witness be a false witness, and hath testified falsely against his brother;

uenerint falsum testem dixisse contra fratrem suum menda-

Liber Deuteronomii

cium 19:19 reddent ei sicut fratri suo facere cogitauit et au- 19:19 Then shall ye do unto him, as he had thought to have

done unto his brother: so shalt thou put the evil away from among you.

feres malum de medio tui 19:20 ut audientes ceteri timorem

19:20 And those which remain shall hear, and fear, and shall henceforth commit no more any such evil among you.

habeant et nequaquam talia audeant facere 19:21 non mise-

19:21 And thine eye shall not pity; but life shall go for life, eye for eye, tooth for tooth, hand for hand, foot for foot.

reberis eius sed animam pro anima oculum pro oculo dentem pro dente manum pro manu pedem pro pede exiges

20:1 si exieris ad bellum contra hostes tuos et uideris equi- 20:1 When thou goest out to battle against thine enemies,

and seest horses, and chariots, and a people more than thou, be not afraid of them: for the LORD thy God is with

tatum et currus et maiorem quam tu habes aduersarii ex- thee, which brought thee up out of the land of Egypt. ercitus multitudinem non timebis eos quia Dominus Deus tuus tecum est qui eduxit te de terra Aegypti 20:2 adpro- 20:2 And it shall be, when ye are come nigh unto the bat-

tle, that the priest shall approach and speak unto the people,

pinquante autem iam proelio stabit sacerdos ante aciem et sic loquetur ad populum 20:3 audi Israhel uos hodie contra in- 20:3 And shall say unto them, Hear, O Israel, ye approach

this day unto battle against your enemies: let not your hearts faint, fear not, and do not tremble, neither be ye

imicos uestros pugnam committitis non pertimescat cor ue- terrified because of them; strum nolite metuere nolite cedere nec formidetis eos 20:4 20:4 For the LORD your God is he that goeth with you, to

fight for you against your enemies, to save you.

quia Dominus Deus uester in medio uestri est et pro uobis contra aduersarios dimicabit ut eruat uos de periculo 20:5 20:5 And the officers shall speak unto the people, saying,

What man is there that hath built a new house, and hath not dedicated it? let him go and return to his house, lest he

duces quoque per singulas turmas audiente exercitu procla- die in the battle, and another man dedicate it. mabunt quis est homo qui aedificauit domum nouam et non dedicauit eam uadat et reuertatur in domum suam ne forte moriatur in bello et alius dedicet illam 20:6 quis est homo 20:6 And what man is he that hath planted a vineyard, and

hath not yet eaten of it? let him also go and return unto his house, lest he die in the battle, and another man eat of it.

qui plantauit uineam et necdum eam fecit esse communem et de qua uesci omnibus liceat uadat et reuertatur in domum suam ne forte moriatur in bello et alius homo eius fungatur officio 20:7 quis est homo qui despondit uxorem et non ac- 20:7 And what man is there that hath betrothed a wife, and

hath not taken her? let him go and return unto his house, lest he die in the battle, and another man take her.

cepit eam uadat et reuertatur in domum suam ne forte moriatur in bello et alius homo accipiat eam 20:8 his dictis addent 20:8 And the officers shall speak further unto the people,

and they shall say, What man is there that is fearful and fainthearted? let him go and return unto his house, lest his

reliqua et loquentur ad populum quis est homo formidolosus brethren’s heart faint as well as his heart. et corde pauido uadat et reuertatur in domum suam ne pauere faciat corda fratrum suorum sicut ipse timore perterritus est 20:9 cumque siluerint exercitus duces et finem loquendi 20:9 And it shall be, when the officers have made an end

of speaking unto the people that they shall make captains of the armies to lead the people.

fecerint unusquisque suos ad bellandum cuneos praeparabit 20:10 si quando accesseris ad expugnandam ciuitatem offeres 20:10 When thou comest nigh unto a city to fight against

it, then proclaim peace unto it.

ei primum pacem 20:11 si receperit et aperuerit tibi portas 20:11 And it shall be, if it make thee answer of peace, and

open unto thee, then it shall be, that all the people that is found therein shall be tributaries unto thee, and they shall

cunctus populus qui in ea est saluabitur et seruiet tibi sub serve thee.

Biblia Sacra Vulgata

20:12 And if it will make no peace with thee, but will make tributo 20:12 sin autem foedus inire noluerint et receperint

war against thee, then thou shalt besiege it:

20:13 And when the LORD thy God hath delivered it into contra te bellum obpugnabis eam 20:13 cumque tradiderit

thine hands, thou shalt smite every male thereof with the edge of the sword:

Dominus Deus tuus illam in manu tua percuties omne quod

20:14 But the women, and the little ones, and the cattle, in ea generis masculini est in ore gladii 20:14 absque mu-

and all that is in the city, even all the spoil thereof, shalt thou take unto thyself; and thou shalt eat the spoil of thine

enemies, which the LORD thy God hath given thee. lieribus et infantibus iumentis et ceteris quae in ciuitate sunt

omnem praedam exercitui diuides et comedes de spoliis ho-

20:15 Thus shalt thou do unto all the cities which are very stium tuorum quae Dominus Deus tuus dederit tibi 20:15 far off from thee, which are not of the cities of these nations.

sic facies cunctis ciuitatibus quae a te procul ualde sunt et non sunt de his urbibus quas in possessionem accepturus es

20:16 But of the cities of these people, which the LORD 20:16 de his autem ciuitatibus quae dabuntur tibi nullum om-

thy God doth give thee for an inheritance, thou shalt save alive nothing that breatheth:

nino permittes uiuere 20:17 sed interficies in ore gladii Het-

20:17 But thou shalt utterly destroy them; namely, the Hittites, and the Amorites, the Canaanites, and the Perizzites, the Hivites, and the Jebusites; as the LORD thy God hath commanded thee:

theum uidelicet et Amorreum et Chananeum Ferezeum et Eueum et Iebuseum sicut praecepit tibi Dominus Deus tuus

20:18 That they teach you not to do after all their abomi- 20:18 ne forte doceant uos facere cunctas abominationes quas

nations, which they have done unto their gods; so should ye sin against the LORD your God.

ipsi operati sunt diis suis et peccetis in Dominum Deum ue-

20:19 When thou shalt besiege a city a long time, in mak- strum 20:19 quando obsederis ciuitatem multo tempore et

ing war against it to take it, thou shalt not destroy the trees thereof by forcing an axe against them: for thou mayest eat of them, and thou shalt not cut them down (for the tree of the field is man’s life) to employ them in the siege:

munitionibus circumdederis ut expugnes eam non succides arbores de quibus uesci potest nec securibus per circuitum debes uastare regionem quoniam lignum est et non homo

20:20 Only the trees which thou knowest that they be not nec potest bellantium contra te augere numerum 20:20 si

trees for meat, thou shalt destroy and cut them down; and thou shalt build bulwarks against the city that maketh war

with thee, until it be subdued. qua autem ligna non sunt pomifera sed agrestia et in ceteros apta usus succide et extrue machinas donec capias ciuitatem quae contra te dimicat

21:1 If one be found slain in the land which the LORD thy 21:1 quando inuentum fuerit in terra quam Dominus De-God giveth thee to possess it, lying in the field, and it be not known who hath slain him:

us tuus daturus est tibi hominis cadauer occisi et ignoratur

21:2 Then thy elders and thy judges shall come forth, and caedis reus 21:2 egredientur maiores natu et iudices tui et

they shall measure unto the cities which are round about him that is slain:

metientur a loco cadaueris singularum per circuitum spatia

21:3 And it shall be, that the city which is next unto the ciuitatum 21:3 et quam uiciniorem ceteris esse perspexe-

slain man, even the elders of that city shall take an heifer, which hath not been wrought with, and which hath not

drawn in the yoke; rint seniores ciuitatis eius tollent uitulam de armento quae

21:4 And the elders of that city shall bring down the heifer non traxit iugum nec terram scidit uomere 21:4 et ducent unto a rough valley, which is neither eared nor sown, and shall strike off the heifer’s neck there in the valley:

eam ad uallem asperam atque saxosam quae numquam arata est nec sementem recepit et caedent in ea ceruices uitulae

21:5 And the priests the sons of Levi shall come near; for 21:5 accedentque sacerdotes filii Leui quos elegerit Dominus

them the LORD thy God hath chosen to minister unto him, and to bless in the name of the LORD; and by their word shall every controversy and every stroke be tried:

Liber Deuteronomii

Deus tuus ut ministrent ei et benedicant in nomine eius et ad uerbum eorum omne negotium et quicquid mundum uel inmundum est iudicetur 21:6 et maiores natu ciuitatis illius 21:6 And all the elders of that city, that are next unto the

slain man, shall wash their hands over the heifer that is beheaded in the valley:

ad interfectum lauabuntque manus suas super uitulam quae in ualle percussa est 21:7 et dicent manus nostrae non ef- 21:7 And they shall answer and say, Our hands have not

shed this blood, neither have our eyes seen it.

fuderunt hunc sanguinem nec oculi uiderunt 21:8 propitius 21:8 Be merciful, O LORD, unto thy people Israel, whom

thou hast redeemed, and lay not innocent blood unto thy people of Israel’s charge. And the blood shall be forgiven

esto populo tuo Israhel quem redeeisti Domine et non re- them. putes sanguinem innocentem in medio populi tui Israhel et auferetur ab eis reatus sanguinis 21:9 tu autem alienus eris 21:9 So shalt thou put away the guilt of innocent blood

from among you, when thou shalt do that which is right in the sight of the LORD.

ab innocentis cruore qui fusus est cum feceris quod praecepit Dominus 21:10 si egressus fueris ad pugnam contra inimi- 21:10 When thou goest forth to war against thine enemies,

and the LORD thy God hath delivered them into thine hands, and thou hast taken them captive,

cos tuos et tradiderit eos Dominus Deus tuus in manu tua captiuosque duxeris 21:11 et uideris in numero captiuorum 21:11 And seest among the captives a beautiful woman,

and hast a desire unto her, that thou wouldest have her to thy wife;

mulierem pulchram et adamaueris eam uoluerisque habere

uxorem 21:12 introduces in domum tuam quae radet cae- 21:12 Then thou shalt bring her home to thine house, and

she shall shave her head, and pare her nails;

sariem et circumcidet ungues 21:13 et deponet uestem in 21:13 And she shall put the raiment of her captivity from

off her, and shall remain in thine house, and bewail her father and her mother a full month: and after that thou shalt go in unto her, and be her husband, and she shall be thy wife.

qua capta est sedensque in domo tua flebit patrem et matrem suam uno mense et postea intrabis ad eam dormiesque cum illa et erit uxor tua 21:14 sin autem postea non sederit 21:14 And it shall be, if thou have no delight in her, then

thou shalt let her go whither she will; but thou shalt not sell her at all for money, thou shalt not make merchandise

animo tuo dimittes eam liberam nec uendere poteris pecunia of her, because thou hast humbled her. nec opprimere per potentiam quia humiliasti eam 21:15 si 21:15 If a man have two wives, one beloved, and another

hated, and they have born him children, both the beloved and the hated; and if the firstborn son be hers that was

habuerit homo uxores duas unam dilectam et alteram odio- hated: sam genuerintque ex eo liberos et fuerit filius odiosae primogenitus 21:16 uolueritque substantiam inter filios suos diui- 21:16 Then it shall be, when he maketh his sons to inherit

that which he hath, that he may not make the son of the beloved firstborn before the son of the hated, which is in-

dere non poterit filium dilectae facere primogenitum et prae- deed the firstborn: ferre filio odiosae 21:17 sed filium odiosae agnoscet primo- 21:17 But he shall acknowledge the son of the hated for

the firstborn, by giving him a double portion of all that he hath: for he is the beginning of his strength; the right of

genitum dabitque ei de his quae habuerit cuncta duplicia iste the firstborn is his. est enim principium liberorum eius et huic debentur primogenita 21:18 si genuerit homo filium contumacem et proter- 21:18 If a man have a stubborn and rebellious son, which

will not obey the voice of his father, or the voice of his mother, and that, when they have chastened him, will not

uum qui non audiat patris aut matris imperium et coercitus hearken unto them: oboedire contempserit 21:19 adprehendent eum et ducent 21:19 Then shall his father and his mother lay hold on him,

and bring him out unto the elders of his city, and unto the gate of his place;

ad seniores ciuitatis illius et ad portam iudicii 21:20 dicent-

21:20 And they shall say unto the elders of his city, This our son is stubborn and rebellious, he will not obey our voice; he is a glutton, and a drunkard.

que ad eos filius noster iste proteruus et contumax est mo-

Biblia Sacra Vulgata

nita nostra audire contemnit comesationibus uacat et luxu-

21:21 And all the men of his city shall stone him with riae atque conuiuiis 21:21 lapidibus eum obruet populus ci-

stones, that he die: so shalt thou put evil away from among you; and all Israel shall hear, and fear.

uitatis et morietur ut auferatis malum de medio uestri et uni-

21:22 And if a man have committed a sin worthy of death, uersus Israhel audiens pertimescat 21:22 quando peccauerit

and he be to be put to death, and thou hang him on a tree:

homo quod morte plectendum est et adiudicatus morti ad-

21:23 His body shall not remain all night upon the tree, pensus fuerit in patibulo 21:23 non permanebit cadauer eius

but thou shalt in any wise bury him that day; (for he that is hanged is accursed of God;) that thy land be not defiled,

which the LORD thy God giveth thee for an inheritance. in ligno sed in eadem die sepelietur quia maledictus a Deo est qui pendet in ligno et nequaquam contaminabis terram tuam quam Dominus Deus tuus dederit tibi in possessionem

22:1 Thou shalt not see thy brother’s ox or his sheep go 22:1 non uidebis bouem fratris tui aut ouem errantem et

astray, and hide thyself from them: thou shalt in any case bring them again unto thy brother.

praeteribis sed reduces fratri tuo 22:2 etiam si non est pro-

22:2 And if thy brother be not nigh unto thee, or if thou know him not, then thou shalt bring it unto thine own house, and it shall be with thee until thy brother seek after it, and thou shalt restore it to him again.

pinquus tuus frater nec nosti eum duces in domum tuam et

22:3 In like manner shalt thou do with his ass; and so shalt erunt apud te quamdiu quaerat ea frater tuus et recipiat 22:3

thou do with his raiment; and with all lost thing of thy brother’s, which he hath lost, and thou hast found, shalt

thou do likewise: thou mayest not hide thyself. similiter facies de asino et de uestimento et de omni re frat-

ris tui quae perierit si inueneris eam ne neglegas quasi alie-

22:4 Thou shalt not see thy brother’s ass or his ox fall nam 22:4 si uideris asinum fratris tui aut bouem cecidisse

down by the way, and hide thyself from them: thou shalt surely help him to lift them up again.

in uia non despicies sed subleuabis cum eo 22:5 non in-

22:5 The woman shall not wear that which pertaineth unto a man, neither shall a man put on a woman’s garment: for all that do so are abomination unto the LORD thy God.

duetur mulier ueste uirili nec uir utetur ueste feminea abo-

22:6 If a bird’s nest chance to be before thee in the way in minabilis enim apud Deum est qui facit haec 22:6 si amany tree, or on the ground, whether they be young ones, or eggs, and the dam sitting upon the young, or upon the

eggs, thou shalt not take the dam with the young: bulans per uiam in arbore uel in terra nidum auis inueneris

et matrem pullis uel ouis desuper incubantem non tenebis

22:7 But thou shalt in any wise let the dam go, and take eam cum filiis 22:7 sed abire patieris captos tenens filios ut

the young to thee; that it may be well with thee, and that thou mayest prolong thy days.

bene sit tibi et longo uiuas tempore 22:8 cum aedificaue-

22:8 When thou buildest a new house, then thou shalt make a battlement for thy roof, that thou bring not blood upon thine house, if any man fall from thence.

ris domum nouam facies murum tecti per circuitum ne effundatur sanguis in domo tua et sis reus labente alio et in

22:9 Thou shalt not sow thy vineyard with divers seeds: praeceps ruente 22:9 non seres uineam tuam altero semine

lest the fruit of thy seed which thou hast sown, and the fruit of thy vineyard, be defiled.

ne et sementis quam seuisti et quae nascuntur ex uinea pa-

22:10 Thou shalt not plow with an ox and an ass together. riter sanctificentur 22:10 non arabis in boue simul et asino 22:11 Thou shalt not wear a garment of divers sorts, as of 22:11 non indueris uestimento quod ex lana linoque contex-

woollen and linen together.

22:12 Thou shalt make thee fringes upon the four quarters tum est 22:12 funiculos in fimbriis facies per quattuor an-

of thy vesture, wherewith thou coverest thyself.

22:13 If any man take a wife, and go in unto her, and hate gulos pallii tui quo operieris 22:13 si duxerit uir uxorem

her,

22:14 And give occasions of speech against her, and bring et postea eam odio habuerit 22:14 quaesieritque occasiones

up an evil name upon her, and say, I took this woman, and when I came to her, I found her not a maid:

quibus dimittat eam obiciens ei nomen pessimum et dixerit

Liber Deuteronomii

uxorem hanc accepi et ingressus ad eam non inueni uirginem 22:15 tollent eam pater et mater eius et ferent secum signa 22:15 Then shall the father of the damsel, and her mother,

take and bring forth the tokens of the damsel’s virginity unto the elders of the city in the gate:

uirginitatis eius ad seniores urbis qui in porta sunt 22:16 et

22:16 And the damsel’s father shall say unto the elders, I gave my daughter unto this man to wife, and he hateth her;

dicet pater filiam meam dedi huic uxorem quam quia odit 22:17 inponet ei nomen pessimum ut dicat non inueni filiam 22:17 And, lo, he hath given occasions of speech against

her, saying, I found not thy daughter a maid; and yet these are the tokens of my daughter’s virginity. And they shall

tuam uirginem et ecce haec sunt signa uirginitatis filiae meae spread the cloth before the elders of the city. expandent uestimentum coram senibus ciuitatis 22:18 ad- 22:18 And the elders of that city shall take that man and

chastise him;

prehendentque senes urbis illius uirum et uerberabunt illum 22:19 condemnantes insuper centum siclis argenti quos da- 22:19 And they shall amerce him in an hundred shekels of

silver, and give them unto the father of the damsel, because he hath brought up an evil name upon a virgin of Israel: and she shall be his wife; he may not put her away all his days.

bit patri puellae quoniam diffamauit nomen pessimum super uirginem Israhel habebitque eam uxorem et non poterit di-

mittere omni tempore uitae suae 22:20 quod si uerum est 22:20 But if this thing be true, and the tokens of virginity

be not found for the damsel:

quod obicit et non est in puella inuenta uirginitas 22:21 ei- 22:21 Then they shall bring out the damsel to the door of

her father’s house, and the men of her city shall stone her with stones that she die: because she hath wrought folly in Israel, to play the whore in her father’s house: so shalt thou put evil away from among you.

cient eam extra fores domus patris sui et lapidibus obruent uiri ciuitatis eius et morietur quoniam fecit nefas in Israhel ut fornicaretur in domo patris sui et auferes malum de medio tui 22:22 si dormierit uir cum uxore alterius uterque mo- 22:22 If a man be found lying with a woman married to

an husband, then they shall both of them die, both the man that lay with the woman, and the woman: so shalt thou put

rientur id est adulter et adultera et auferes malum de Isra- away evil from Israel. hel 22:23 si puellam uirginem desponderit uir et inuenerit 22:23 If a damsel that is a virgin be betrothed unto an hus-

band, and a man find her in the city, and lie with her;

eam aliquis in ciuitate et concubuerit cum illa 22:24 edu- 22:24 Then ye shall bring them both out unto the gate of

that city, and ye shall stone them with stones that they die; the damsel, because she cried not, being in the city; and the man, because he hath humbled his neighbour’s wife: so thou shalt put away evil from among you.

ces utrumque ad portam ciuitatis illius et lapidibus obruentur puella quia non clamauit cum esset in ciuitate uir quia humiliauit uxorem proximi sui et auferes malum de medio tui 22:25 sin autem in agro reppererit uir puellam quae despon- 22:25 But if a man find a betrothed damsel in the field, and

the man force her, and lie with her: then the man only that lay with her shall die.

sata est et adprehendens concubuerit cum illa ipse morietur solus 22:26 puella nihil patietur nec est rea mortis quoniam 22:26 But unto the damsel thou shalt do nothing; there is

in the damsel no sin worthy of death: for as when a man riseth against his neighbour, and slayeth him, even so is

sicut latro consurgit contra fratrem suum et occidit animam this matter: eius ita et puella perpessa est 22:27 sola erat in agro clama- 22:27 For he found her in the field, and the betrothed

damsel cried, and there was none to save her.

uit et nullus adfuit qui liberaret eam 22:28 si inuenerit uir 22:28 If a man find a damsel that is a virgin, which is not

betrothed, and lay hold on her, and lie with her, and they be found;

puellam uirginem quae non habet sponsum et adprehendens concubuerit cum ea et res ad iudicium uenerit 22:29 dabit 22:29 Then the man that lay with her shall give unto the

damsel’s father fifty shekels of silver, and she shall be his wife; because he hath humbled her, he may not put her

qui dormiuit cum ea patri puellae quinquaginta siclos argenti away all his days. et habebit eam uxorem quia humiliauit illam non poterit di-

Biblia Sacra Vulgata

22:30 A man shall not take his father’s wife, nor discover mittere cunctis diebus uitae suae 22:30 non accipiet homo

his father’s skirt.

uxorem patris sui nec reuelabit operimentum eius

23:1 He that is wounded in the stones, or hath his privy 23:1 non intrabit eunuchus adtritis uel amputatis testicu-

member cut off, shall not enter into the congregation of the LORD.

lis et absciso ueretro ecclesiam Domini 23:2 non ingredie-

23:2 A bastard shall not enter into the congregation of the LORD; even to his tenth generation shall he not enter into the congregation of the LORD.

tur mamzer hoc est de scorto natus in ecclesiam Domini us-

23:3 An Ammonite or Moabite shall not enter into the con- que ad decimam generationem 23:3 Ammanites et Moabigregation of the LORD; even to their tenth generation shall they not enter into the congregation of the LORD for ever:

tes etiam post decimam generationem non intrabunt eccle-

23:4 Because they met you not with bread and with water siam Domini in aeternum 23:4 quia noluerunt uobis occur-

in the way, when ye came forth out of Egypt; and because they hired against thee Balaam the son of Beor of Pethor

of Mesopotamia, to curse thee. rere cum pane et aqua in uia quando egressi estis de Aegypto

et quia conduxerunt contra te Balaam filium Beor de Meso-

23:5 Nevertheless the LORD thy God would not hearken potamiam Syriae ut malediceret tibi 23:5 et noluit Domi-

unto Balaam; but the LORD thy God turned the curse into a blessing unto thee, because the LORD thy God loved

thee. nus Deus tuus audire Balaam uertitque maledictionem eius 23:6 Thou shalt not seek their peace nor their prosperity in benedictionem tuam eo quod diligeret te 23:6 non facies

all thy days for ever.

cum eis pacem nec quaeres eis bona cunctis diebus uitae tuae

23:7 Thou shalt not abhor an Edomite; for he is thy in sempiternum 23:7 non abominaberis Idumeum quia fra-

brother: thou shalt not abhor an Egyptian; because thou wast a stranger in his land.

ter tuus est nec Aegyptium quia aduena fuisti in terra eius

23:8 The children that are begotten of them shall enter into 23:8 qui nati fuerint ex eis tertia generatione intrabunt eccle-

the congregation of the LORD in their third generation. 23:9 When the host goeth forth against thine enemies, then siam Domini 23:9 quando egressus fueris aduersus hostes keep thee from every wicked thing.

23:10 If there be among you any man, that is not clean by tuos in pugnam custodies te ab omni re mala 23:10 si fue-

reason of uncleanness that chanceth him by night, then shall he go abroad out of the camp, he shall not come

within the camp: rit inter uos homo qui nocturno pollutus sit somnio egredie-23:11 But it shall be, when evening cometh on, he shall tur extra castra 23:11 et non reuertetur priusquam ad ues-

wash himself with water: and when the sun is down, he shall come into the camp again.

peram lauetur aqua et post solis occasum regredietur in cas-

23:12 Thou shalt have a place also without the camp, tra 23:12 habebis locum extra castra ad quem egrediaris ad

whither thou shalt go forth abroad:

23:13 And thou shalt have a paddle upon thy weapon; and requisita naturae 23:13 gerens paxillum in balteo cumque

it shall be, when thou wilt ease thyself abroad, thou shalt dig therewith, and shalt turn back and cover that which

cometh from thee: sederis fodies per circuitum et egesta humo operies 23:14

23:14 For the LORD thy God walketh in the midst of thy camp, to deliver thee, and to give up thine enemies before thee; therefore shall thy camp be holy: that he see no unclean thing in thee, and turn away from thee.

quo releuatus es Dominus enim Deus tuus ambulat in medio castrorum ut eruat te et tradat tibi inimicos tuos ut sint castra tua sancta et nihil in eis appareat foeditatis nec de-

23:15 Thou shalt not deliver unto his master the servant relinquat te 23:15 non trades seruum domino suo qui ad te which is escaped from his master unto thee:

23:16 He shall dwell with thee, even among you, in that confugerit 23:16 habitabit tecum in loco qui ei placuerit et

place which he shall choose in one of thy gates, where it liketh him best: thou shalt not oppress him.

in una urbium tuarum requiescet nec contristes eum 23:17

23:17 There shall be no whore of the daughters of Israel, nor a sodomite of the sons of Israel.

non erit meretrix de filiabus Israhel neque scortator de filiis

23:18 Thou shalt not bring the hire of a whore, or the price Israhel 23:18 non offeres mercedem prostibuli nec pretium

of a dog, into the house of the LORD thy God for any vow: for even both these are abomination unto the LORD thy God.

Liber Deuteronomii

canis in domum Domini Dei tui quicquid illud est quod uouerint quia abominatio est utrumque apud Dominum Deum

tuum 23:19 non fenerabis fratri tuo ad usuram pecuniam 23:19 Thou shalt not lend upon usury to thy brother; usury of money, usury of victuals, usury of any thing that is lent upon usury:

nec fruges nec quamlibet aliam rem 23:20 sed alieno fratri

23:20 Unto a stranger thou mayest lend upon usury; but unto thy brother thou shalt not lend upon usury: that the LORD thy God may bless thee in all that thou settest thine hand to in the land whither thou goest to possess it.

autem tuo absque usura id quod indiget commodabis ut benedicat tibi Dominus Deus tuus in omni opere tuo in terra ad quam ingredieris possidendam 23:21 cum uoueris uo- 23:21 When thou shalt vow a vow unto the LORD thy

God, thou shalt not slack to pay it: for the LORD thy God will surely require it of thee; and it would be sin in thee.

tum Domino Deo tuo non tardabis reddere quia requiret illud Dominus Deus tuus et si moratus fueris reputabit tibi in peccatum 23:22 si nolueris polliceri absque peccato eris 23:23 23:22 But if thou shalt forbear to vow, it shall be no sin in

thee.

23:23 That which is gone out of thy lips thou shalt keep and perform; even a freewill offering, according as thou hast vowed unto the LORD thy God, which thou hast promised with thy mouth.

quod autem semel egressum est de labiis tuis obseruabis et facies sicut promisisti Domino Deo tuo et propria uoluntate et ore tuo locutus es 23:24 ingressus uineam proximi tui co- 23:24 When thou comest into thy neighbour’s vineyard,

then thou mayest eat grapes thy fill at thine own pleasure; but thou shalt not put any in thy vessel.

mede uuas quantum tibi placuerit foras autem ne efferas tecum 23:25 si intraueris in segetem amici tui franges spicas 23:25 When thou comest into the standing corn of thy

neighbour, then thou mayest pluck the ears with thine hand; but thou shalt not move a sickle unto thy neighbour’s

et manu conteres falce autem non metes standing corn.

24:1 si acceperit homo uxorem et habuerit eam et non in- 24:1 When a man hath taken a wife, and married her, and

it come to pass that she find no favour in his eyes, because he hath found some uncleanness in her: then let him write her a bill of divorcement, and give it in her hand, and send her out of his house.

uenerit gratiam ante oculos eius propter aliquam foeditatem scribet libellum repudii et dabit in manu illius et dimittet eam de domo sua 24:2 cumque egressa alterum maritum duxerit 24:2 And when she is departed out of his house, she may

go and be another man’s wife.

24:3 et ille quoque oderit eam dederitque ei libellum repudii 24:3 And if the latter husband hate her, and write her a bill

of divorcement, and giveth it in her hand, and sendeth her out of his house; or if the latter husband die, which took

et dimiserit de domo sua uel certe mortuus fuerit 24:4 non her to be his wife;

24:4 Her former husband, which sent her away, may not take her again to be his wife, after that she is defiled; for that is abomination before the LORD: and thou shalt not cause the land to sin, which the LORD thy God giveth thee for an inheritance.

poterit prior maritus recipere eam in uxorem quia polluta est et abominabilis facta est coram Domino ne peccare facias terram tuam quam Dominus Deus tuus tibi tradiderit possidendam 24:5 cum acceperit homo nuper uxorem non pro- 24:5 When a man hath taken a new wife, he shall not go

out to war, neither shall he be charged with any business: but he shall be free at home one year, and shall cheer up

cedet ad bellum nec ei quippiam necessitatis iniungetur pub- his wife which he hath taken. licae sed uacabit absque culpa domui suae ut uno anno laetetur cum uxore sua 24:6 non accipies loco pignoris inferio- 24:6 No man shall take the nether or the upper millstone

to pledge: for he taketh a man’s life to pledge.

rem et superiorem molam quia animam suam adposuit tibi 24:7 si deprehensus fuerit homo sollicitans fratrem suum de 24:7 If a man be found stealing any of his brethren of the

children of Israel, and maketh merchandise of him, or selleth him; then that thief shall die; and thou shalt put evil

filiis Israhel et uendito eo accipiens pretium interficietur et away from among you. auferes malum de medio tui 24:8 obserua diligenter ne in- 24:8 Take heed in the plague of leprosy, that thou ob-

serve diligently, and do according to all that the priests the Levites shall teach you: as I commanded them, so ye shall observe to do.

Biblia Sacra Vulgata

curras in plagam leprae sed facies quaecumque docuerint te sacerdotes leuitici generis iuxta id quod praecepi eis et imple

24:9 Remember what the LORD thy God did unto Miriam sollicite 24:9 mementote quae fecerit Dominus Deus uester

by the way, after that ye were come forth out of Egypt.

24:10 When thou dost lend thy brother any thing, thou Mariae in uia cum egrederemini de Aegypto 24:10 cum re-

shalt not go into his house to fetch his pledge.

petes a proximo tuo rem aliquam quam debet tibi non ingre-

24:11 Thou shalt stand abroad, and the man to whom thou dieris domum eius ut pignus auferas 24:11 sed stabis foris

dost lend shall bring out the pledge abroad unto thee.

24:12 And if the man be poor, thou shalt not sleep with his et ille tibi proferet quod habuerit 24:12 sin autem pauper

pledge:

24:13 In any case thou shalt deliver him the pledge again est non pernoctabit apud te pignus 24:13 sed statim reddes

when the sun goeth down, that he may sleep in his own raiment, and bless thee: and it shall be righteousness unto

thee before the LORD thy God. ei ante solis occasum ut dormiens in uestimento suo bene-24:14 Thou shalt not oppress an hired servant that is dicat tibi et habeas iustitiam coram Domino Deo tuo 24:14

poor and needy, whether he be of thy brethren, or of thy strangers that are in thy land within thy gates:

non negabis mercedem indigentis et pauperis fratris tui siue aduenae qui tecum moratur in terra et intra portas tuas est

24:15 At his day thou shalt give him his hire, neither shall 24:15 sed eadem die reddes ei pretium laboris sui ante solis

the sun go down upon it; for he is poor, and setteth his heart upon it: lest he cry against thee unto the LORD, and

it be sin unto thee. occasum quia pauper est et ex eo sustentat animam suam ne

clamet contra te ad Dominum et reputetur tibi in peccatum

24:16 The fathers shall not be put to death for the children, 24:16 non occidentur patres pro filiis nec filii pro patribus sed

neither shall the children be put to death for the fathers: every man shall be put to death for his own sin.

unusquisque pro suo peccato morietur 24:17 non peruertes

24:17 Thou shalt not pervert the judgment of the stranger, nor of the fatherless; nor take a widow’s raiment to pledge:

iudicium aduenae et pupilli nec auferes pignoris loco uiduae

24:18 But thou shalt remember that thou wast a bondman uestimentum 24:18 memento quod seruieris in Aegypto et

in Egypt, and the LORD thy God redeemed thee thence: therefore I command thee to do this thing.

eruerit te Dominus Deus tuus inde idcirco praecipio tibi ut

24:19 When thou cuttest down thine harvest in thy field, facias hanc rem 24:19 quando messueris segetem in agro and hast forgot a sheaf in the field, thou shalt not go again to fetch it: it shall be for the stranger, for the fatherless, and for the widow: that the LORD thy God may bless thee in all the work of thine hands.

tuo et oblitus manipulum reliqueris non reuerteris ut tollas eum sed aduenam et pupillum et uiduam auferre patieris ut benedicat tibi Dominus Deus tuus in omni opere manuum

24:20 When thou beatest thine olive tree, thou shalt not go tuarum 24:20 si fruges colliges oliuarum quicquid remanse-

over the boughs again: it shall be for the stranger, for the fatherless, and for the widow.

rit in arboribus non reuerteris ut colligas sed relinques adue-

24:21 When thou gatherest the grapes of thy vineyard, nae pupillo ac uiduae 24:21 si uindemiaueris uineam tuam

thou shalt not glean it afterward: it shall be for the stranger, for the fatherless, and for the widow.

non colliges remanentes racemos sed cedent in usus adue-

24:22 And thou shalt remember that thou wast a bondman nae pupilli ac uiduae 24:22 memento quod et tu seruieris in

in the land of Egypt: therefore I command thee to do this thing.

Aegypto et idcirco praecipiam tibi ut facias hanc rem

25:1 If there be a controversy between men, and they come 25:1 si fuerit causa inter aliquos et interpellauerint iudi-

unto judgment, that the judges may judge them; then they shall justify the righteous, and condemn the wicked.

ces quem iustum esse perspexerint illi iustitiae palmam da-

25:2 And it shall be, if the wicked man be worthy to be bunt quem impium condemnabunt impietatis 25:2 sin au-

beaten, that the judge shall cause him to lie down, and to be beaten before his face, according to his fault, by a certain number.

Liber Deuteronomii

tem eum qui peccauit dignum uiderint plagis prosternent et coram se facient uerberari pro mensura peccati erit et pla-

garum modus 25:3 ita dumtaxat ut quadragenarium nume- 25:3 Forty stripes he may give him, and not exceed: lest, if he should exceed, and beat him above these with many stripes, then thy brother should seem vile unto thee.

rum non excedant ne foede laceratus ante oculos tuos abeat frater tuus 25:4 non ligabis os bouis terentis in area fru- 25:4 Thou shalt not muzzle the ox when he treadeth out

the corn.

ges tuas 25:5 quando habitauerint fratres simul et unus ex 25:5 If brethren dwell together, and one of them die, and

have no child, the wife of the dead shall not marry without unto a stranger: her husband’s brother shall go in unto her, and take her to him to wife, and perform the duty of an husband’s brother unto her.

eis absque liberis mortuus fuerit uxor defuncti non nubet alteri sed accipiet eam frater eius et suscitabit semen fratris sui 25:6 et primogenitum ex ea filium nomine illius appel- 25:6 And it shall be, that the firstborn which she beareth

shall succeed in the name of his brother which is dead, that his name be not put out of Israel.

labit ut non deleatur nomen eius ex Israhel 25:7 sin autem

25:7 And if the man like not to take his brother’s wife, then let his brother’s wife go up to the gate unto the elders, and say, My husband’s brother refuseth to raise up unto his brother a name in Israel, he will not perform the duty of my husband’s brother.

noluerit accipere uxorem fratris sui quae ei lege debetur perget mulier ad portam ciuitatis et interpellabit maiores natu dicetque non uult frater uiri mei suscitare nomen fratris sui in Israhel nec me in coniugium sumere 25:8 statimque ac- 25:8 Then the elders of his city shall call him, and speak

unto him: and if he stand to it, and say, I like not to take her;

cersiri eum facient et interrogabunt si responderit nolo eam uxorem accipere 25:9 accedet mulier ad eum coram senio- 25:9 Then shall his brother’s wife come unto him in the

presence of the elders, and loose his shoe from off his foot, and spit in his face, and shall answer and say, So shall it be done unto that man that will not build up his brother’s house.

ribus et tollet calciamentum de pede eius spuetque in faciem illius et dicet sic fit homini qui non aedificat domum fratris sui 25:10 et uocabitur nomen illius in Israhel domus Discul- 25:10 And his name shall be called in Israel, The house of

him that hath his shoe loosed.

ciati 25:11 si habuerint inter se iurgium uiri et unus contra 25:11 When men strive together one with another, and the

wife of the one draweth near for to deliver her husband out of the hand of him that smiteth him, and putteth forth her

alterum rixari coeperit uolensque uxor alterius eruere uirum hand, and taketh him by the secrets: suum de manu fortioris miserit manum et adprehenderit uerenda eius 25:12 abscides manum illius nec flecteris super 25:12 Then thou shalt cut off her hand, thine eye shall not

pity her.

eam ulla misericordia 25:13 non habebis in sacculo diuersa 25:13 Thou shalt not have in thy bag divers weights, a

great and a small.

pondera maius et minus 25:14 nec erit in domo tua modius 25:14 Thou shalt not have in thine house divers measures,

a great and a small.

maior et minor 25:15 pondus habebis iustum et uerum et 25:15 But thou shalt have a perfect and just weight, a perfect and just measure shalt thou have: that thy days may be lengthened in the land which the LORD thy God giveth

modius aequalis et uerus erit tibi ut multo uiuas tempore su- thee. per terram quam Dominus Deus tuus dederit tibi 25:16 abo- 25:16 For all that do such things, and all that do unrigh-

teously, are an abomination unto the LORD thy God.

minatur enim Dominus eum qui facit haec et auersatur omnem iniustitiam 25:17 memento quae fecerit tibi Amalech 25:17 Remember what Amalek did unto thee by the way,

when ye were come forth out of Egypt;

in uia quando egrediebaris ex Aegypto 25:18 quomodo oc- 25:18 How he met thee by the way, and smote the hind-

most of thee, even all that were feeble behind thee, when thou wast faint and weary; and he feared not God.

currerit tibi et extremos agminis tui qui lassi residebant ceciderit quando tu eras fame et labore confectus et non timuerit

Biblia Sacra Vulgata

25:19 Therefore it shall be, when the LORD thy God hath Deum 25:19 cum ergo Dominus Deus tuus dederit tibi regiven thee rest from all thine enemies round about, in the land which the LORD thy God giveth thee for an inheritance to possess it, that thou shalt blot out the remembrance of Amalek from under heaven; thou shalt not forget it.

quiem et subiecerit cunctas per circuitum nationes in terra quam tibi pollicitus est delebis nomen eius sub caelo caue ne obliuiscaris

26:1 And it shall be, when thou art come in unto the land 26:1 cumque intraueris terram quam Dominus Deus tuus which the LORD thy God giveth thee for an inheritance, and possessest it, and dwellest therein;

tibi daturus est possidendam et obtinueris eam atque habita-

26:2 That thou shalt take of the first of all the fruit of the ueris in illa 26:2 tolles de cunctis frugibus primitias et po-

earth, which thou shalt bring of thy land that the LORD thy God giveth thee, and shalt put it in a basket, and shalt go unto the place which the LORD thy God shall choose to place his name there.

nes in cartallo pergesque ad locum quem Dominus Deus tuus

26:3 And thou shalt go unto the priest that shall be in those elegerit ut ibi inuocetur nomen eius 26:3 accedesque ad sa-

days, and say unto him, I profess this day unto the LORD thy God, that I am come unto the country which the LORD

sware unto our fathers for to give us. cerdotem qui fuerit in diebus illis et dices ad eum profiteor

hodie coram Domino Deo tuo quod ingressus sim terram pro

26:4 And the priest shall take the basket out of thine hand, qua iurauit patribus nostris ut daret eam nobis 26:4 susci-

and set it down before the altar of the LORD thy God.

piensque sacerdos cartallum de manu eius ponet ante altare

26:5 And thou shalt speak and say before the LORD thy Domini Dei tui 26:5 et loqueris in conspectu Domini Dei

God, A Syrian ready to perish was my father, and he went down into Egypt, and sojourned there with a few, and be-

came there a nation, great, mighty, and populous: tui Syrus persequebatur patrem meum qui descendit in Aegyptum et ibi peregrinatus est in paucissimo numero creuitque in gentem magnam et robustam et infinitae multitudinis

26:6 And the Egyptians evil entreated us, and afflicted us, 26:6 adflixeruntque nos Aegyptii et persecuti sunt inponentes

and laid upon us hard bondage:

26:7 And when we cried unto the LORD God of our fa- onera grauissima 26:7 et clamauimus ad Dominum Deum

thers, the LORD heard our voice, and looked on our affliction, and our labour, and our oppression:

patrum nostrorum qui exaudiuit nos et respexit humilitatem

26:8 And the LORD brought us forth out of Egypt with a nostram et laborem atque angustias 26:8 et eduxit nos de

mighty hand, and with an outstretched arm, and with great terribleness, and with signs, and with wonders:

Aegypto in manu forti et brachio extento in ingenti pauore

26:9 And he hath brought us into this place, and hath given in signis atque portentis 26:9 et introduxit ad locum istum

us this land, even a land that floweth with milk and honey.

26:10 And now, behold, I have brought the firstfruits of et tradidit nobis terram lacte et melle manantem 26:10 et

the land, which thou, O LORD, hast given me. And thou shalt set it before the LORD thy God, and worship before

the LORD thy God: idcirco nunc offero primitias frugum terrae quam dedit Do-

minus mihi et dimittes eas in conspectu Domini Dei tui ad-

26:11 And thou shalt rejoice in every good thing which orato Domino Deo tuo 26:11 et epulaberis in omnibus bo-

the LORD thy God hath given unto thee, and unto thine house, thou, and the Levite, and the stranger that is among

you. nis quae Dominus Deus tuus dederit tibi et domui tuae tu et 26:12 When thou hast made an end of tithing all the tithes Leuites et aduena qui tecum est 26:12 quando conpleueris

of thine increase the third year, which is the year of tithing, and hast given it unto the Levite, the stranger, the fatherless, and the widow, that they may eat within thy gates, and be filled;

decimam cunctarum frugum tuarum anno decimarum tertio dabis Leuitae et aduenae et pupillo et uiduae ut comedant in-

26:13 Then thou shalt say before the LORD thy God, I tra portas tuas et saturentur 26:13 loquerisque in conspectu

have brought away the hallowed things out of mine house, and also have given them unto the Levite, and unto the stranger, to the fatherless, and to the widow, according to all thy commandments which thou hast commanded me: I have not transgressed thy commandments, neither have I forgotten them.

Domini Dei tui abstuli quod sanctificatum est de domo mea

Liber Deuteronomii

et dedi illud Leuitae et aduenae pupillo et uiduae sicut iussisti mihi non praeteriui mandata tua nec sum oblitus imperii 26:14 non comedi ex eis in luctu meo nec separaui ea 26:14 I have not eaten thereof in my mourning, neither

have I taken away ought thereof for any unclean use, nor given ought thereof for the dead: but I have hearkened to the voice of the LORD my God, and have done according to all that thou hast commanded me.

in qualibet inmunditia nec expendi ex his quicquam in re funebri oboediui uoci Domini Dei mei et feci omnia sicut praecepisti mihi 26:15 respice de sanctuario tuo de excelso 26:15 Look down from thy holy habitation, from heaven,

and bless thy people Israel, and the land which thou hast given us, as thou swarest unto our fathers, a land that

caelorum habitaculo et benedic populo tuo Israhel et terrae floweth with milk and honey. quam dedisti nobis sicut iurasti patribus nostris terrae lacte et melle mananti 26:16 hodie Dominus Deus tuus praece- 26:16 This day the LORD thy God hath commanded thee

to do these statutes and judgments: thou shalt therefore keep and do them with all thine heart, and with all thy

pit tibi ut facias mandata haec atque iudicia et custodias et soul. impleas ex toto corde tuo et ex tota anima tua 26:17 Domi- 26:17 Thou hast avouched the LORD this day to be thy

God, and to walk in his ways, and to keep his statutes, and his commandments, and his judgments, and to hear-

num elegisti hodie ut sit tibi Deus et ambules in uiis eius et ken unto his voice: custodias caerimonias illius et mandata atque iudicia et oboedias eius imperio 26:18 et Dominus elegit te hodie ut sis 26:18 And the LORD hath avouched thee this day to be his

peculiar people, as he hath promised thee, and that thou shouldest keep all his commandments;

ei populus peculiaris sicut locutus est tibi et custodias omnia praecepta eius 26:19 et faciat te excelsiorem cunctis genti- 26:19 And to make thee high above all nations which he

hath made, in praise, and in name, and in honour; and that thou mayest be an holy people unto the LORD thy God,

bus quas creauit in laudem et nomen et gloriam suam ut sis as he hath spoken. populus sanctus Domini Dei tui sicut locutus est

27:1 praecepit autem Moses et seniores Israhel populo di- 27:1 And Moses with the elders of Israel commanded the

people, saying, Keep all the commandments which I command you this day.

centes custodite omne mandatum quod praecipio uobis hodie 27:2 cumque transieritis Iordanem in terram quam Do- 27:2 And it shall be on the day when ye shall pass over

Jordan unto the land which the LORD thy God giveth thee, that thou shalt set thee up great stones, and plaister them

minus Deus tuus dabit tibi eriges ingentes lapides et calce le- with plaister: uigabis eos 27:3 ut possis in eis scribere omnia uerba legis 27:3 And thou shalt write upon them all the words of this

law, when thou art passed over, that thou mayest go in unto the land which the LORD thy God giveth thee, a land that floweth with milk and honey; as the LORD God of thy fathers hath promised thee.

huius Iordane transmisso ut introeas terram quam Dominus Deus tuus dabit tibi terram lacte et melle manantem sicut iurauit patribus tuis 27:4 quando ergo transieritis Iordanem 27:4 Therefore it shall be when ye be gone over Jordan,

that ye shall set up these stones, which I command you this day, in mount Ebal, and thou shalt plaister them with

erige lapides quos ego hodie praecipio uobis in monte Hebal plaister. et leuigabis calce 27:5 et aedificabis ibi altare Domino Deo 27:5 And there shalt thou build an altar unto the LORD

thy God, an altar of stones: thou shalt not lift up any iron tool upon them.

tuo de lapidibus quos ferrum non tetigit 27:6 et de saxis

27:6 Thou shalt build the altar of the LORD thy God of whole stones: and thou shalt offer burnt offerings thereon unto the LORD thy God:

informibus et inpolitis et offeres super eo holocausta Domino Deo tuo 27:7 et immolabis hostias pacificas comedes- 27:7 And thou shalt offer peace offerings, and shalt eat

there, and rejoice before the LORD thy God.

que ibi et epulaberis coram Domino Deo tuo 27:8 et scribes 27:8 And thou shalt write upon the stones all the words of

this law very plainly.

super lapides omnia uerba legis huius plane et lucide 27:9 27:9 And Moses and the priests the Levites spake unto all

Israel, saying, Take heed, and hearken, O Israel; this day thou art become the people of the LORD thy God.

Biblia Sacra Vulgata

dixeruntque Moses et sacerdotes leuitici generis ad omnem Israhelem adtende et audi Israhel hodie factus es populus

27:10 Thou shalt therefore obey the voice of the LORD Domini Dei tui 27:10 audies uocem eius et facies mandata

thy God, and do his commandments and his statutes, which I command thee this day.

atque iustitias quas ego praecipio tibi 27:11 praecepitque

27:11 And Moses charged the people the same day, saying,

27:12 These shall stand upon mount Gerizim to bless the Moses populo in die illo dicens 27:12 hii stabunt ad benedi-

people, when ye are come over Jordan; Simeon, and Levi, and Judah, and Issachar, and Joseph, and Benjamin:

cendum Domino super montem Garizim Iordane transmisso

27:13 And these shall stand upon mount Ebal to curse; Symeon Leui Iudas Isachar Ioseph et Beniamin 27:13 et e

Reuben, Gad, and Asher, and Zebulun, Dan, and Naphtali.

regione isti stabunt ad maledicendum in monte Hebal Ruben

27:14 And the Levites shall speak, and say unto all the Gad et Aser Zabulon Dan et Nepthalim 27:14 et pronuntia-

men of Israel with a loud voice,

bunt Leuitae dicentque ad omnes uiros Israhel excelsa uoce

27:15 Cursed be the man that maketh any graven or molten 27:15 maledictus homo qui facit sculptile et conflatile abo-

image, an abomination unto the LORD, the work of the hands of the craftsman, and putteth it in a secret place.

And all the people shall answer and say, Amen. minationem Domini opus manuum artificum ponetque illud

in abscondito et respondebit omnis populus et dicet amen

27:16 Cursed be he that setteth light by his father or his 27:16 maledictus qui non honorat patrem suum et matrem mother. And all the people shall say, Amen.

27:17 Cursed be he that removeth his neighbour’s land- et dicet omnis populus amen 27:17 maledictus qui transfert

mark. And all the people shall say, Amen. 27:18 Cursed be he that maketh the blind to wander out of terminos proximi sui et dicet omnis populus amen 27:18 the way. And all the people shall say, Amen.

maledictus qui errare facit caecum in itinere et dicet omnis

27:19 Cursed be he that perverteth the judgment of the populus amen 27:19 maledictus qui peruertit iudicium ad-

stranger, fatherless, and widow. And all the people shall say, Amen.

uenae pupilli et uiduae et dicet omnis populus amen 27:20

27:20 Cursed be he that lieth with his father’s wife; because he uncovereth his father’s skirt. And all the people shall say, Amen.

maledictus qui dormit cum uxore patris sui et reuelat ope-

27:21 Cursed be he that lieth with any manner of beast. rimentum lectuli eius et dicet omnis populus amen 27:21

And all the people shall say, Amen.

maledictus qui dormit cum omni iumento et dicet omnis po-

27:22 Cursed be he that lieth with his sister, the daughter pulus amen 27:22 maledictus qui dormit cum sorore sua fiof his father, or the daughter of his mother. And all the people shall say, Amen.

lia patris sui siue matris suae et dicet omnis populus amen

27:23 Cursed be he that lieth with his mother in law. And 27:23 maledictus qui dormit cum socru sua et dicet omnis

all the people shall say, Amen.

27:24 Cursed be he that smiteth his neighbour secretly. populus amen 27:24 maledictus qui clam percusserit proxi-

And all the people shall say, Amen. 27:25 Cursed be he that taketh reward to slay an innocent mum suum et dicet omnis populus amen 27:25 maledictus person. And all the people shall say, Amen.

qui accipit munera ut percutiat animam sanguinis innocen-

27:26 Cursed be he that confirmeth not all the words of tis et dicet omnis populus amen 27:26 maledictus qui non

this law to do them. And all the people shall say, Amen.

permanet in sermonibus legis huius nec eos opere perficit et dicet omnis populus amen

28:1 And it shall come to pass, if thou shalt hearken dili- 28:1 sin autem audieris uocem Domini Dei tui ut facias gently unto the voice of the LORD thy God, to observe and to do all his commandments which I command thee this day, that the LORD thy God will set thee on high above all nations of the earth:

atque custodias omnia mandata eius quae ego praecipio tibi

Liber Deuteronomii

hodie faciet te Dominus Deus tuus excelsiorem cunctis gentibus quae uersantur in terra 28:2 uenientque super te uni- 28:2 And all these blessings shall come on thee, and over-

take thee, if thou shalt hearken unto the voice of the LORD thy God.

uersae benedictiones istae et adprehendent te si tamen praecepta eius audieris 28:3 benedictus tu in ciuitate et bene- 28:3 Blessed shalt thou be in the city, and blessed shalt

thou be in the field.

dictus in agro 28:4 benedictus fructus uentris tui et fructus 28:4 Blessed shall be the fruit of thy body, and the fruit of

thy ground, and the fruit of thy cattle, the increase of thy kine, and the flocks of thy sheep.

terrae tuae fructusque iumentorum tuorum greges armentorum et caulae ouium tuarum 28:5 benedicta horrea tua et 28:5 Blessed shall be thy basket and thy store. benedictae reliquiae tuae 28:6 benedictus eris et ingrediens 28:6 Blessed shalt thou be when thou comest in, and

blessed shalt thou be when thou goest out.

et egrediens 28:7 dabit Dominus inimicos tuos qui consur- 28:7 The LORD shall cause thine enemies that rise up

against thee to be smitten before thy face: they shall come out against thee one way, and flee before thee seven ways.

gunt aduersum te corruentes in conspectu tuo per unam uiam uenient contra te et per septem fugient a facie tua 28:8 emit- 28:8 The LORD shall command the blessing upon thee

in thy storehouses, and in all that thou settest thine hand unto; and he shall bless thee in the land which the LORD

tet Dominus benedictionem super cellaria tua et super om- thy God giveth thee. nia opera manuum tuarum benedicetque tibi in terra quam acceperis 28:9 suscitabit te Dominus sibi in populum sanc- 28:9 The LORD shall establish thee an holy people unto

himself, as he hath sworn unto thee, if thou shalt keep the commandments of the LORD thy God, and walk in his

tum sicut iurauit tibi si custodieris mandata Domini Dei tui ways. et ambulaueris in uiis eius 28:10 uidebuntque omnes terra- 28:10 And all people of the earth shall see that thou art

called by the name of the LORD; and they shall be afraid of thee.

rum populi quod nomen Domini inuocatum sit super te et timebunt te 28:11 abundare te faciet Dominus omnibus bo- 28:11 And the LORD shall make thee plenteous in goods,

in the fruit of thy body, and in the fruit of thy cattle, and in the fruit of thy ground, in the land which the LORD sware

nis fructu uteri tui et fructu iumentorum tuorum fructu terrae unto thy fathers to give thee. tuae quam iurauit Dominus patribus tuis ut daret tibi 28:12 28:12 The LORD shall open unto thee his good treasure,

the heaven to give the rain unto thy land in his season, and to bless all the work of thine hand: and thou shalt lend

aperiet Dominus thesaurum suum optimum caelum ut tribuat unto many nations, and thou shalt not borrow. pluuiam terrae tuae in tempore suo benedicet cunctis operibus manuum tuarum et fenerabis gentibus multis et ipse a nullo fenus accipies 28:13 constituet te Dominus in caput 28:13 And the LORD shall make thee the head, and not

the tail; and thou shalt be above only, and thou shalt not be beneath; if that thou hearken unto the commandments of the LORD thy God, which I command thee this day, to observe and to do them:

et non in caudam et eris semper supra et non subter si audieris mandata Domini Dei tui quae ego praecipio tibi hodie et custodieris et feceris 28:14 ac non declinaueris ab eis 28:14 And thou shalt not go aside from any of the words

which I command thee this day, to the right hand, or to the left, to go after other gods to serve them.

nec ad dextram nec ad sinistram nec secutus fueris deos alienos neque colueris eos 28:15 quod si audire nolueris uocem 28:15 But it shall come to pass, if thou wilt not hearken

unto the voice of the LORD thy God, to observe to do all his commandments and his statutes which I command thee this day; that all these curses shall come upon thee, and overtake thee:

Domini Dei tui ut custodias et facias omnia mandata eius et caerimonias quas ego praecipio tibi hodie uenient super te omnes maledictiones istae et adprehendent te 28:16 male- 28:16 Cursed shalt thou be in the city, and cursed shalt

thou be in the field.

dictus eris in ciuitate maledictus in agro 28:17 maledictum 28:17 Cursed shall be thy basket and thy store.

Biblia Sacra Vulgata

28:18 Cursed shall be the fruit of thy body, and the fruit horreum tuum et maledictae reliquiae tuae 28:18 maledic-

of thy land, the increase of thy kine, and the flocks of thy sheep.

tus fructus uentris tui et fructus terrae tuae armenta boum

28:19 Cursed shalt thou be when thou comest in, and tuorum et greges ouium tuarum 28:19 maledictus eris in-

cursed shalt thou be when thou goest out.

28:20 The LORD shall send upon thee cursing, vexation, grediens et maledictus egrediens 28:20 mittet Dominus su-

and rebuke, in all that thou settest thine hand unto for to do, until thou be destroyed, and until thou perish quickly; because of the wickedness of thy doings, whereby thou hast forsaken me.

per te famem et esuriem et increpationem in omnia opera tua quae facies donec conterat te et perdat uelociter propter ad-

28:21 The LORD shall make the pestilence cleave unto inuentiones tuas pessimas in quibus reliquisti me 28:21 ad-

thee, until he have consumed thee from off the land, whither thou goest to possess it.

iungat Dominus tibi pestilentiam donec consumat te de terra

28:22 The LORD shall smite thee with a consumption, and ad quam ingredieris possidendam 28:22 percutiat te Domi-

with a fever, and with an inflammation, and with an extreme burning, and with the sword, and with blasting, and

with mildew; and they shall pursue thee until thou perish. nus egestate febri et frigore ardore et aestu et aere corrupto 28:23 And thy heaven that is over thy head shall be brass, ac robigine et persequatur donec pereas 28:23 sit caelum

and the earth that is under thee shall be iron.

28:24 The LORD shall make the rain of thy land powder quod supra te est aeneum et terra quam calcas ferrea 28:24

and dust: from heaven shall it come down upon thee, until thou be destroyed.

det Dominus imbrem terrae tuae puluerem et de caelo de-

28:25 The LORD shall cause thee to be smitten before scendat super te cinis donec conteraris 28:25 tradat te Do-

thine enemies: thou shalt go out one way against them, and flee seven ways before them: and shalt be removed

into all the kingdoms of the earth. minus corruentem ante hostes tuos per unam uiam egrediaris

contra eos et per septem fugias et dispergaris per omnia re-

28:26 And thy carcase shall be meat unto all fowls of the gna terrae 28:26 sitque cadauer tuum in escam cunctis uola-

air, and unto the beasts of the earth, and no man shall fray them away.

tilibus caeli et bestiis terrae et non sit qui abigat 28:27 per-

28:27 The LORD will smite thee with the botch of Egypt, and with the emerods, and with the scab, and with the itch, whereof thou canst not be healed.

cutiat te Dominus ulcere Aegypti et parte corporis per quam stercora digeruntur scabie quoque et prurigine ita ut curari

28:28 The LORD shall smite thee with madness, and nequeas 28:28 percutiat te Dominus amentia et caecitate ac

blindness, and astonishment of heart:

28:29 And thou shalt grope at noonday, as the blind furore mentis 28:29 et palpes in meridie sicut palpare solet

gropeth in darkness, and thou shalt not prosper in thy ways: and thou shalt be only oppressed and spoiled ev-

ermore, and no man shall save thee. caecus in tenebris et non dirigas uias tuas omnique tempore

calumniam sustineas et opprimaris uiolentia nec habeas qui

28:30 Thou shalt betroth a wife, and another man shall liberet te 28:30 uxorem accipias et alius dormiat cum ea

lie with her: thou shalt build an house, and thou shalt not dwell therein: thou shalt plant a vineyard, and shalt not

gather the grapes thereof. domum aedifices et non habites in ea plantes uineam et non

28:31 Thine ox shall be slain before thine eyes, and thou uindemies eam 28:31 bos tuus immoletur coram te et non shalt not eat thereof: thine ass shall be violently taken away from before thy face, and shall not be restored to thee: thy sheep shall be given unto thine enemies, and thou shalt have none to rescue them.

comedas ex eo asinus tuus rapiatur in conspectu tuo et non reddatur tibi oues tuae dentur inimicis tuis et non sit qui te

28:32 Thy sons and thy daughters shall be given unto an- adiuuet 28:32 filii tui et filiae tuae tradantur alteri populo

other people, and thine eyes shall look, and fail with longing for them all the day long; and there shall be no might

in thine hand. uidentibus oculis tuis et deficientibus ad conspectum eorum 28:33 The fruit of thy land, and all thy labours, shall a tota die et non sit fortitudo in manu tua 28:33 fructus ter-

nation which thou knowest not eat up; and thou shalt be only oppressed and crushed alway:

rae tuae et omnes labores tuos comedat populus quem ig-

Liber Deuteronomii

noras et sis semper calumniam sustinens et oppressus cunctis diebus 28:34 et stupens ad terrorem eorum quae uide- 28:34 So that thou shalt be mad for the sight of thine eyes

which thou shalt see.

bunt oculi tui 28:35 percutiat te Dominus ulcere pessimo 28:35 The LORD shall smite thee in the knees, and in the

legs, with a sore botch that cannot be healed, from the sole of thy foot unto the top of thy head.

in genibus et in suris sanarique non possis a planta pedis

usque ad uerticem tuum 28:36 ducet Dominus te et regem 28:36 The LORD shall bring thee, and thy king which thou shalt set over thee, unto a nation which neither thou nor thy fathers have known; and there shalt thou serve other gods,

tuum quem constitueris super te in gentem quam ignoras tu wood and stone. et patres tui et seruies ibi diis alienis ligno et lapidi 28:37 28:37 And thou shalt become an astonishment, a proverb,

and a byword, among all nations whither the LORD shall lead thee.

et eris perditus in prouerbium ac fabulam omnibus populis ad quos te introduxerit Dominus 28:38 sementem multam 28:38 Thou shalt carry much seed out into the field, and

shalt gather but little in; for the locust shall consume it.

iacies in terram et modicum congregabis quia lucustae om-

nia deuorabunt 28:39 uineam plantabis et fodies et uinum 28:39 Thou shalt plant vineyards, and dress them, but shalt neither drink of the wine, nor gather the grapes; for the worms shall eat them.

non bibes nec colliges ex ea quippiam quoniam uastabitur

uermibus 28:40 oliuas habebis in omnibus terminis tuis et 28:40 Thou shalt have olive trees throughout all thy coasts, but thou shalt not anoint thyself with the oil; for thine olive shall cast his fruit.

non ungueris oleo quia defluent et peribunt 28:41 filios ge-

28:41 Thou shalt beget sons and daughters, but thou shalt not enjoy them; for they shall go into captivity.

nerabis et filias et non frueris eis quoniam ducentur in captiuitatem 28:42 omnes arbores tuas et fruges terrae tuae 28:42 All thy trees and fruit of thy land shall the locust

consume.

robigo consumet 28:43 aduena qui tecum uersatur in terra 28:43 The stranger that is within thee shall get up above

thee very high; and thou shalt come down very low.

ascendet super te eritque sublimior tu autem descendes et eris inferior 28:44 ipse fenerabit tibi et tu non fenerabis ei 28:44 He shall lend to thee, and thou shalt not lend to him:

he shall be the head, and thou shalt be the tail.

ipse erit in caput et tu eris in caudam 28:45 et uenient su- 28:45 Moreover all these curses shall come upon thee,

and shall pursue thee, and overtake thee, till thou be destroyed; because thou hearkenedst not unto the voice of the LORD thy God, to keep his commandments and his statutes which he commanded thee:

per te omnes maledictiones istae et persequentes adprehendent te donec intereas quia non audisti uocem Domini Dei tui nec seruasti mandata eius et caerimonias quas praecepit tibi 28:46 et erunt in te signa atque prodiiia et in semine tuo 28:46 And they shall be upon thee for a sign and for a

wonder, and upon thy seed for ever.

usque in sempiternum 28:47 eo quod non seruieris Domino 28:47 Because thou servedst not the LORD thy God with

joyfulness, and with gladness of heart, for the abundance of all things;

Deo tuo in gaudio cordisque laetitia propter rerum omnium abundantiam 28:48 seruies inimico tuo quem inmittet Do- 28:48 Therefore shalt thou serve thine enemies which the

LORD shall send against thee, in hunger, and in thirst, and in nakedness, and in want of all things: and he shall put a

minus tibi in fame et siti et nuditate et omnium penuria et yoke of iron upon thy neck, until he have destroyed thee. ponet iugum ferreum super ceruicem tuam donec te conterat 28:49 adducet Dominus super te gentem de longinquo 28:49 The LORD shall bring a nation against thee from

far, from the end of the earth, as swift as the eagle flieth; a nation whose tongue thou shalt not understand;

et de extremis finibus terrae in similitudinem aquilae uolantis cum impetu cuius linguam intellegere non possis 28:50 28:50 A nation of fierce countenance, which shall not re-

gard the person of the old, nor shew favour to the young:

gentem procacissimam quae non deferat seni nec miserea-

Biblia Sacra Vulgata

28:51 And he shall eat the fruit of thy cattle, and the fruit tur paruulo 28:51 et deuoret fructum iumentorum tuorum of thy land, until thou be destroyed: which also shall not leave thee either corn, wine, or oil, or the increase of thy

kine, or flocks of thy sheep, until he have destroyed thee. ac fruges terrae tuae donec intereas et non relinquat tibi tri-

ticum uinum et oleum armenta boum et greges ouium donec

28:52 And he shall besiege thee in all thy gates, until thy te disperdat 28:52 et conterat in cunctis urbibus tuis et de-

high and fenced walls come down, wherein thou trustedst, throughout all thy land: and he shall besiege thee in all thy gates throughout all thy land, which the LORD thy God hath given thee.

struantur muri tui firmi atque sublimes in quibus habebas fiduciam in omni terra tua obsideberis intra portas tuas in

28:53 And thou shalt eat the fruit of thine own body, the omni terra quam dabit tibi Dominus Deus tuus 28:53 et co-

flesh of thy sons and of thy daughters, which the LORD thy God hath given thee, in the siege, and in the straitness,

wherewith thine enemies shall distress thee: medes fructum uteri tui et carnes filiorum et filiarum tuarum

quas dedit tibi Dominus Deus tuus in angustia et uastitate

28:54 So that the man that is tender among you, and very qua opprimet te hostis tuus 28:54 homo delicatus in te et

delicate, his eye shall be evil toward his brother, and toward the wife of his bosom, and toward the remnant of his

children which he shall leave: luxuriosus ualde inuidebit fratri suo et uxori quae cubat in 28:55 So that he will not give to any of them of the flesh of sinu suo 28:55 ne det eis de carnibus filiorum suorum quas

his children whom he shall eat: because he hath nothing left him in the siege, and in the straitness, wherewith thine

enemies shall distress thee in all thy gates. comedet eo quod nihil habeat aliud in obsidione et penuria 28:56 The tender and delicate woman among you, which qua uastauerint te inimici tui intra omnes portas tuas 28:56

would not adventure to set the sole of her foot upon the ground for delicateness and tenderness, her eye shall be evil toward the husband of her bosom, and toward her son, and toward her daughter,

tenera mulier et delicata quae super terram ingredi non ualebat nec pedis uestigium figere propter mollitiem et teneritudinem nimiam inuidebit uiro suo qui cubat in sinu eius su-

28:57 And toward her young one that cometh out from per filii et filiae carnibus 28:57 et inluuie secundarum quae

between her feet, and toward her children which she shall bear: for she shall eat them for want of all things secretly in the siege and straitness, wherewith thine enemy shall distress thee in thy gates.

egrediuntur de medio feminum eius et super liberis qui eadem hora nati sunt comedent enim eos clam propter rerum omnium penuriam in obsidione et uastitate qua opprimet te

28:58 If thou wilt not observe to do all the words of this inimicus tuus intra portas tuas 28:58 nisi custodieris et fe-

law that are written in this book, that thou mayest fear this glorious and fearful name, THE LORD THY GOD;

ceris omnia uerba legis huius quae scripta sunt in hoc uolumine et timueris nomen eius gloriosum et terribile hoc est

28:59 Then the LORD will make thy plagues wonderful, Dominum Deum tuum 28:59 augebit Dominus plagas tuas and the plagues of thy seed, even great plagues, and of long continuance, and sore sicknesses, and of long contin-

uance. et plagas seminis tui plagas magnas et perseuerantes infirmi-28:60 Moreover he will bring upon thee all the diseases tates pessimas et perpetuas 28:60 et conuertet in te omnes

of Egypt, which thou wast afraid of; and they shall cleave unto thee.

adflictiones Aegypti quas timuisti et adherebunt tibi 28:61

28:61 Also every sickness, and every plague, which is not written in the book of this law, them will the LORD bring upon thee, until thou be destroyed.

insuper et uniuersos languores et plagas quae non sunt scriptae in uolumine legis huius inducet Dominus super te do-

28:62 And ye shall be left few in number, whereas ye were nec te conterat 28:62 et remanebitis pauci numero qui prius

as the stars of heaven for multitude; because thou wouldest not obey the voice of the LORD thy God.

eratis sicut astra caeli prae multitudine quoniam non audisti

28:63 And it shall come to pass, that as the LORD re- uocem Domini Dei tui 28:63 et sicut ante laetatus est Do-

joiced over you to do you good, and to multiply you; so the LORD will rejoice over you to destroy you, and to bring you to nought; and ye shall be plucked from off the land whither thou goest to possess it.

Liber Deuteronomii

minus super uos bene uobis faciens uosque multiplicans sic laetabitur disperdens uos atque subuertens ut auferamini de terra ad quam ingredieris possidendam 28:64 disperget te 28:64 And the LORD shall scatter thee among all people,

from the one end of the earth even unto the other; and there thou shalt serve other gods, which neither thou nor

Dominus in omnes populos a summitate terrae usque ad ter- thy fathers have known, even wood and stone. minos eius et seruies ibi diis alienis quos et tu ignoras et patres tui lignis et lapidibus 28:65 in gentibus quoque illis non 28:65 And among these nations shalt thou find no ease,

neither shall the sole of thy foot have rest: but the LORD shall give thee there a trembling heart, and failing of eyes,

quiesces neque erit requies uestigio pedis tui dabit enim tibi and sorrow of mind: Dominus ibi cor pauidum et deficientes oculos et animam maerore consumptam 28:66 et erit uita tua quasi pendens 28:66 And thy life shall hang in doubt before thee; and

thou shalt fear day and night, and shalt have none assurance of thy life:

ante te timebis nocte et die et non credes uitae tuae 28:67

28:67 In the morning thou shalt say, Would God it were even! and at even thou shalt say, Would God it were morning! for the fear of thine heart wherewith thou shalt fear, and for the sight of thine eyes which thou shalt see.

mane dices quis mihi det uesperum et uespere quis mihi det mane propter cordis tui formidinem qua terreberis et propter ea quae tuis uidebis oculis 28:68 reducet te Dominus clas- 28:68 And the LORD shall bring thee into Egypt again

with ships, by the way whereof I spake unto thee, Thou shalt see it no more again: and there ye shall be sold unto your enemies for bondmen and bondwomen, and no man shall buy you.

sibus in Aegyptum per uiam de qua dixi tibi ut eam amplius non uideres ibi uenderis inimicis tuis in seruos et ancillas et non erit qui emat

29:1 haec sunt uerba foederis quod praecepit Dominus 29:1 These are the words of the covenant, which the

LORD commanded Moses to make with the children of Israel in the land of Moab, beside the covenant which he

Mosi ut feriret cum filiis Israhel in terra Moab praeter il- made with them in Horeb.

lud foedus quod cum eis pepigit in Horeb 29:2 uocauitque 29:2 And Moses called unto all Israel, and said unto them, Ye have seen all that the LORD did before your eyes in the land of Egypt unto Pharaoh, and unto all his servants, and

Moses omnem Israhelem et dixit ad eos uos uidistis uniuersa unto all his land; quae fecit Dominus coram uobis in terra Aegypti Pharaoni et omnibus seruis eius uniuersaeque terrae illius 29:3 tem- 29:3 The great temptations which thine eyes have seen, the

signs, and those great miracles:

ptationes magnas quas uiderunt oculi tui signa illa portentaque ingentia 29:4 et non dedit Dominus uobis cor intelle- 29:4 Yet the LORD hath not given you an heart to per-

ceive, and eyes to see, and ears to hear, unto this day.

gens et oculos uidentes et aures quae possint audire usque in praesentem diem 29:5 adduxi uos quadraginta annis per 29:5 And I have led you forty years in the wilderness: your

clothes are not waxen old upon you, and thy shoe is not waxen old upon thy foot.

desertum non sunt adtrita uestimenta uestra nec calciamenta

pedum tuorum uetustate consumpta sunt 29:6 panem non 29:6 Ye have not eaten bread, neither have ye drunk wine or strong drink: that ye might know that I am the LORD your God.

comedistis uinum et siceram non bibistis ut sciretis quia ego sum Dominus Deus uester 29:7 et uenistis ad locum hunc 29:7 And when ye came unto this place, Sihon the king of

Heshbon, and Og the king of Bashan, came out against us unto battle, and we smote them:

egressusque est Seon rex Esebon et Og rex Basan occurrens nobis ad pugnam et percussimus eos 29:8 et tulimus terram 29:8 And we took their land, and gave it for an inheritance

unto the Reubenites, and to the Gadites, and to the half tribe of Manasseh.

eorum ac tradidimus possidendam Ruben et Gad et dimidiae

Biblia Sacra Vulgata

29:9 Keep therefore the words of this covenant, and do tribui Manasse 29:9 custodite ergo uerba pacti huius et im-

them, that ye may prosper in all that ye do.

29:10 Ye stand this day all of you before the LORD your plete ea ut intellegatis uniuersa quae facitis 29:10 uos statis

God; your captains of your tribes, your elders, and your officers, with all the men of Israel,

hodie cuncti coram Domino Deo uestro principes uestri ac tribus et maiores natu atque doctores omnis populus Israhel

29:11 Your little ones, your wives, and thy stranger that is 29:11 liberi et uxores uestrae et aduena qui tecum moratur in thy camp, from the hewer of thy wood unto the drawer of thy water:

in castris exceptis lignorum caesoribus et his qui conpor-

29:12 That thou shouldest enter into covenant with the tant aquas 29:12 ut transeas in foedere Domini Dei tui et

LORD thy God, and into his oath, which the LORD thy God maketh with thee this day:

in iureiurando quod hodie Dominus Deus tuus percutit te-

29:13 That he may establish thee to day for a people unto cum 29:13 ut suscitet te sibi in populum et ipse sit Deus himself, and that he may be unto thee a God, as he hath said unto thee, and as he hath sworn unto thy fathers, to

Abraham, to Isaac, and to Jacob. tuus sicut locutus est tibi et sicut iurauit patribus tuis Ab-29:14 Neither with you only do I make this covenant and raham Isaac et Iacob 29:14 nec uobis solis ego hoc foedus

this oath;

29:15 But with him that standeth here with us this day ferio et haec iuramenta confirmo 29:15 sed cunctis praesen-

before the LORD our God, and also with him that is not here with us this day:

tibus et absentibus 29:16 uos enim nostis ut habitauerimus

29:16 (For ye know how we have dwelt in the land of Egypt; and how we came through the nations which ye passed by;

in terra Aegypti et quomodo transierimus per medium natio-

29:17 And ye have seen their abominations, and their num quas transeuntes 29:17 uidistis abominationes et sor-

idols, wood and stone, silver and gold, which were among them:)

des id est idola eorum lignum et lapidem argentum et aurum

29:18 Lest there should be among you man, or woman, or quae colebant 29:18 ne forte sit inter uos uir aut mulier fa-

family, or tribe, whose heart turneth away this day from the LORD our God, to go and serve the gods of these nations; lest there should be among you a root that beareth gall and wormwood;

milia aut tribus cuius cor auersum est hodie a Domino Deo uestro ut uadat et seruiat diis illarum gentium et sit inter uos

29:19 And it come to pass, when he heareth the words radix germinans fel et amaritudinem 29:19 cumque audie-

of this curse, that he bless himself in his heart, saying, I shall have peace, though I walk in the imagination of mine

heart, to add drunkenness to thirst: rit uerba iuramenti huius benedicat sibi in corde suo dicens

pax erit mihi et ambulabo in prauitate cordis mei et adsu-

29:20 The LORD will not spare him, but then the anger mat ebria sitientem 29:20 et Dominus non ignoscat ei sed

of the LORD and his jealousy shall smoke against that man, and all the curses that are written in this book shall lie upon him, and the LORD shall blot out his name from under heaven.

tunc quam maxime furor eius fumet et zelus contra hominem illum et sedeant super eo omnia maledicta quae scripta

29:21 And the LORD shall separate him unto evil out of sunt in hoc uolumine et deleat nomen eius sub caelo 29:21

all the tribes of Israel, according to all the curses of the covenant that are written in this book of the law:

et consumat eum in perditionem ex omnibus tribubus Israhel iuxta maledictiones quae in libro legis huius ac foede-

29:22 So that the generation to come of your children that ris continentur 29:22 dicetque sequens generatio et filii qui

shall rise up after you, and the stranger that shall come from a far land, shall say, when they see the plagues of that land, and the sicknesses which the LORD hath laid upon it;

nascentur deinceps et peregrini qui de longe uenerint uidentes plagas terrae illius et infirmitates quibus eam adflixerit

29:23 And that the whole land thereof is brimstone, and Dominus 29:23 sulphure et salis ardore conburens ita ut ul-

salt, and burning, that it is not sown, nor beareth, nor any grass groweth therein, like the overthrow of Sodom, and Gomorrah, Admah, and Zeboim, which the LORD overthrew in his anger, and in his wrath:

tra non seratur nec uirens quippiam germinet in exemplum

Liber Deuteronomii

subuersionis Sodomae et Gomorrae Adamae et Seboim quas subuertit Dominus in ira et furore suo 29:24 et dicent om- 29:24 Even all nations shall say, Wherefore hath the

LORD done thus unto this land? what meaneth the heat of this great anger?

nes gentes quare sic fecit Dominus terrae huic quae est haec ira furoris eius inmensa 29:25 et respondebunt quia dere- 29:25 Then men shall say, Because they have forsaken

the covenant of the LORD God of their fathers, which he made with them when he brought them forth out of the

liquerunt pactum Domini quod pepigit cum patribus eorum land of Egypt: quando eduxit eos de terra Aegypti 29:26 et seruierunt diis 29:26 For they went and served other gods, and wor-

shipped them, gods whom they knew not, and whom he had not given unto them:

alienis et adorauerunt eos quos nesciebant et quibus non fuerant adtributi 29:27 idcirco iratus est furor Domini contra 29:27 And the anger of the LORD was kindled against this

land, to bring upon it all the curses that are written in this book:

terram istam ut induceret super eam omnia maledicta quae in hoc uolumine scripta sunt 29:28 et eiecit eos de terra sua 29:28 And the LORD rooted them out of their land in

anger, and in wrath, and in great indignation, and cast them into another land, as it is this day.

in ira et furore et indignatione maxima proiecitque in terram alienam sicut hodie conprobatur 29:29 abscondita Domino 29:29 The secret things belong unto the LORD our God:

but those things which are revealed belong unto us and to our children for ever, that we may do all the words of this

Deo nostro quae manifesta sunt nobis et filiis nostris usque law. in aeternum ut faciamus uniuersa legis huius

30:1 cum ergo uenerint super te omnes sermones isti be- 30:1 And it shall come to pass, when all these things are

come upon thee, the blessing and the curse, which I have set before thee, and thou shalt call them to mind among all

nedictio siue maledictio quam proposui in conspectu tuo et the nations, whither the LORD thy God hath driven thee, ductus paenitudine cordis tui in uniuersis gentibus in quas disperserit te Dominus Deus tuus 30:2 reuersus fueris ad 30:2 And shalt return unto the LORD thy God, and shalt

obey his voice according to all that I command thee this day, thou and thy children, with all thine heart, and with

eum et oboedieris eius imperiis sicut ego hodie praecipio all thy soul; tibi cum filiis tuis in toto corde tuo et in tota anima tua 30:3 30:3 That then the LORD thy God will turn thy captivity,

and have compassion upon thee, and will return and gather thee from all the nations, whither the LORD thy God hath

reducet Dominus Deus tuus captiuitatem tuam ac miserebi- scattered thee. tur tui et rursum congregabit te de cunctis populis in quos te ante dispersit 30:4 si ad cardines caeli fueris dissipatus 30:4 If any of thine be driven out unto the outmost parts of

heaven, from thence will the LORD thy God gather thee, and from thence will he fetch thee:

inde te retrahet Dominus Deus tuus 30:5 et adsumet atque

30:5 And the LORD thy God will bring thee into the land which thy fathers possessed, and thou shalt possess it; and he will do thee good, and multiply thee above thy fathers.

introducet in terram quam possederunt patres tui et obtinebis eam et benedicens tibi maioris numeri esse te faciet quam fuerunt patres tui 30:6 circumcidet Dominus Deus tuus cor 30:6 And the LORD thy God will circumcise thine heart,

and the heart of thy seed, to love the LORD thy God with all thine heart, and with all thy soul, that thou mayest live.

tuum et cor seminis tui ut diligas Dominum Deum tuum in toto corde tuo et in tota anima tua et possis uiuere 30:7 om- 30:7 And the LORD thy God will put all these curses upon

thine enemies, and on them that hate thee, which persecuted thee.

nes autem maledictiones has conuertet super inimicos tuos et eos qui oderunt te et persequuntur 30:8 tu autem reuerte- 30:8 And thou shalt return and obey the voice of the

LORD, and do all his commandments which I command thee this day.

ris et audies uocem Domini Dei tui faciesque uniuersa mandata quae ego praecipio tibi hodie 30:9 et abundare te faciet 30:9 And the LORD thy God will make thee plenteous in

every work of thine hand, in the fruit of thy body, and in the fruit of thy cattle, and in the fruit of thy land, for good: for the LORD will again rejoice over thee for good, as he rejoiced over thy fathers:

Biblia Sacra Vulgata

Dominus Deus tuus in cunctis operibus manuum tuarum in subole uteri tui et in fructu iumentorum tuorum in ubertate terrae tuae et in rerum omnium largitate reuertetur enim Dominus ut gaudeat super te in omnibus bonis sicut gauisus est

30:10 If thou shalt hearken unto the voice of the LORD thy in patribus tuis 30:10 si tamen audieris uocem Domini Dei

God, to keep his commandments and his statutes which are written in this book of the law, and if thou turn unto the LORD thy God with all thine heart, and with all thy soul.

tui et custodieris praecepta eius et caerimonias quae in hac lege conscriptae sunt et reuertaris ad Dominum Deum tuum

30:11 For this commandment which I command thee this in toto corde tuo et in tota anima tua 30:11 mandatum hoc

day, it is not hidden from thee, neither is it far off.

quod ego praecipio tibi hodie non supra te est neque procul

30:12 It is not in heaven, that thou shouldest say, Who positum 30:12 nec in caelo situm ut possis dicere quis no-

shall go up for us to heaven, and bring it unto us, that we may hear it, and do it?

strum ad caelum ualet conscendere ut deferat illud ad nos et

30:13 Neither is it beyond the sea, that thou shouldest say, audiamus atque opere conpleamus 30:13 neque trans mare Who shall go over the sea for us, and bring it unto us, that we may hear it, and do it?

positum ut causeris et dicas quis e nobis transfretare poterit mare et illud ad nos usque deferre ut possimus audire et fa-

30:14 But the word is very nigh unto thee, in thy mouth, cere quod praeceptum est 30:14 sed iuxta te est sermo ualde

and in thy heart, that thou mayest do it.

30:15 See, I have set before thee this day life and good, in ore tuo et in corde tuo ut facias illum 30:15 considera

and death and evil;

quod hodie proposuerim in conspectu tuo uitam et bonum

30:16 In that I command thee this day to love the LORD et e contrario mortem et malum 30:16 ut diligas Dominum

thy God, to walk in his ways, and to keep his commandments and his statutes and his judgments, that thou mayest live and multiply: and the LORD thy God shall bless thee in the land whither thou goest to possess it.

Deum tuum et ambules in uiis eius et custodias mandata illius et caerimonias atque iudicia et uiuas ac multiplicet te benedicatque tibi in terra ad quam ingredieris possidendam

30:17 But if thine heart turn away, so that thou wilt not 30:17 sin autem auersum fuerit cor tuum et audire nolueris

hear, but shalt be drawn away, and worship other gods, and serve them;

atque errore deceptus adoraueris deos alienos et seruieris eis

30:18 I denounce unto you this day, that ye shall surely 30:18 praedico tibi hodie quod pereas et paruo tempore mo-

perish, and that ye shall not prolong your days upon the land, whither thou passest over Jordan to go to possess it.

reris in terra ad quam Iordane transmisso ingredieris possi-

30:19 I call heaven and earth to record this day against dendam 30:19 testes inuoco hodie caelum et terram quod you, that I have set before you life and death, blessing and cursing: therefore choose life, that both thou and thy seed

may live: proposuerim uobis uitam et mortem bonum et malum bene-

dictionem et maledictionem elige ergo uitam ut et tu uiuas et

30:20 That thou mayest love the LORD thy God, and that semen tuum 30:20 et diligas Dominum Deum tuum atque thou mayest obey his voice, and that thou mayest cleave unto him: for he is thy life, and the length of thy days: that thou mayest dwell in the land which the LORD sware unto thy fathers, to Abraham, to Isaac, and to Jacob, to give them.

oboedias uoci eius et illi adhereas ipse est enim uita tua et longitudo dierum tuorum ut habites in terra pro qua iurauit Dominus patribus tuis Abraham Isaac et Iacob ut daret eam illis

Liber Deuteronomii

31:1 abiit itaque Moses et locutus est omnia uerba haec ad 31:1 And Moses went and spake these words unto all Is-

rael.

uniuersum Israhel 31:2 et dixit ad eos centum uiginti anno- 31:2 And he said unto them, I am an hundred and twenty

years old this day; I can no more go out and come in: also the LORD hath said unto me, Thou shalt not go over this

rum sum hodie non possum ultra egredi et ingredi praesertim Jordan. cum et Dominus dixerit mihi non transibis Iordanem istum 31:3 Dominus ergo Deus tuus transibit ante te ipse delebit 31:3 The LORD thy God, he will go over before thee, and

he will destroy these nations from before thee, and thou shalt possess them: and Joshua, he shall go over before

omnes gentes has in conspectu tuo et possidebis eas et Iosue thee, as the LORD hath said. iste transibit ante te sicut locutus est Dominus 31:4 faciet- 31:4 And the LORD shall do unto them as he did to Sihon

and to Og, kings of the Amorites, and unto the land of them, whom he destroyed.

que Dominus eis sicut fecit Seon et Og regibus Amorreorum et terrae eorum delebitque eos 31:5 cum ergo et hos tradi- 31:5 And the LORD shall give them up before your face,

that ye may do unto them according unto all the commandments which I have commanded you.

derit uobis similiter facietis eis sicut praecepi uobis 31:6 ui-

31:6 Be strong and of a good courage, fear not, nor be afraid of them: for the LORD thy God, he it is that doth go with thee; he will not fail thee, nor forsake thee.

riliter agite et confortamini nolite timere nec paueatis a conspectu eorum quia Dominus Deus tuus ipse est ductor tuus et non dimittet nec derelinquet te 31:7 uocauitque Moses Io- 31:7 And Moses called unto Joshua, and said unto him in

the sight of all Israel, Be strong and of a good courage: for thou must go with this people unto the land which the LORD hath sworn unto their fathers to give them; and thou shalt cause them to inherit it.

sue et dixit ei coram omni Israhel confortare et esto robustus tu enim introduces populum istum in terram quam daturum se patribus eorum iurauit Dominus et tu eam sorte diuides 31:8 et Dominus qui ductor uester est ipse erit tecum non 31:8 And the LORD, he it is that doth go before thee; he

will be with thee, he will not fail thee, neither forsake thee: fear not, neither be dismayed.

dimittet nec derelinquet te noli timere nec paueas 31:9 sc-

31:9 And Moses wrote this law, and delivered it unto the priests the sons of Levi, which bare the ark of the covenant of the LORD, and unto all the elders of Israel.

ripsit itaque Moses legem hanc et tradidit eam sacerdotibus filiis Leui qui portabant arcam foederis Domini et cunctis senioribus Israhelis 31:10 praecepitque eis dicens post sep- 31:10 And Moses commanded them, saying, At the end of

every seven years, in the solemnity of the year of release, in the feast of tabernacles,

tem annos anno remissionis in sollemnitate tabernaculorum 31:11 conuenientibus cunctis ex Israhel ut appareant in con- 31:11 When all Israel is come to appear before the LORD

thy God in the place which he shall choose, thou shalt read this law before all Israel in their hearing.

spectu Domini Dei tui in loco quem elegerit Dominus leges uerba legis huius coram omni Israhel audientibus eis 31:12 31:12 Gather the people together, men and women, and

children, and thy stranger that is within thy gates, that they may hear, and that they may learn, and fear the LORD

et in unum omni populo congregato tam uiris quam mulieri- your God, and observe to do all the words of this law: bus paruulis et aduenis qui sunt intra portas tuas ut audientes discant et timeant Dominum Deum uestrum et custodiant impleantque omnes sermones legis huius 31:13 filii quoque 31:13 And that their children, which have not known any

thing, may hear, and learn to fear the LORD your God, as long as ye live in the land whither ye go over Jordan to

eorum qui nunc ignorant audire possint et timeant Dominum possess it. Deum suum cunctis diebus quibus uersantur in terra ad quam uos Iordane transito pergitis obtinendam 31:14 et ait Domi- 31:14 And the LORD said unto Moses, Behold, thy days

approach that thou must die: call Joshua, and present yourselves in the tabernacle of the congregation, that I may give him a charge. And Moses and Joshua went, and presented themselves in the tabernacle of the congregation.

nus ad Mosen ecce prope sunt dies mortis tuae uoca Iosue

Biblia Sacra Vulgata

et state in tabernaculo testimonii ut praecipiam ei abierunt ergo Moses et Iosue et steterunt in tabernaculo testimonii

31:15 And the LORD appeared in the tabernacle in a pillar 31:15 apparuitque Dominus ibi in columna nubis quae stetit

of a cloud: and the pillar of the cloud stood over the door of the tabernacle.

in introitu tabernaculi 31:16 dixitque Dominus ad Mosen

31:16 And the LORD said unto Moses, Behold, thou shalt sleep with thy fathers; and this people will rise up, and go a whoring after the gods of the strangers of the land, whither they go to be among them, and will forsake me, and break my covenant which I have made with them.

ecce tu dormies cum patribus tuis et populus iste consurgens fornicabitur post deos alienos in terra ad quam ingredietur et habitabit in ea ibi derelinquet me et irritum faciet

31:17 Then my anger shall be kindled against them in that foedus quod pepigi cum eo 31:17 et irascetur furor meus day, and I will forsake them, and I will hide my face from them, and they shall be devoured, and many evils and troubles shall befall them; so that they will say in that day, Are not these evils come upon us, because our God is not among us?

contra eum in die illo et derelinquam eum et abscondam faciem meam ab eo et erit in deuorationem inuenient eum omnia mala et adflictiones ita ut dicat in illo die uere quia non

31:18 And I will surely hide my face in that day for all est Deus mecum inuenerunt me haec mala 31:18 ego authe evils which they shall have wrought, in that they are turned unto other gods.

tem abscondam et celabo faciem meam in die illo propter

31:19 Now therefore write ye this song for you, and teach omnia mala quae fecit quia secutus est deos alienos 31:19

it the children of Israel: put it in their mouths, that this song may be a witness for me against the children of Israel.

nunc itaque scribite uobis canticum istud et docete filios Israhel ut memoriter teneant et ore decantent et sit mihi car-

31:20 For when I shall have brought them into the land men istud pro testimonio inter filios Israhel 31:20 introdu-

which I sware unto their fathers, that floweth with milk and honey; and they shall have eaten and filled themselves, and waxen fat; then will they turn unto other gods, and serve them, and provoke me, and break my covenant.

cam enim eum in terram pro qua iuraui patribus eius lacte et melle manantem cumque comederint et saturati crassique fuerint auertentur ad deos alienos et seruient eis et detrahent

31:21 And it shall come to pass, when many evils and trou- mihi et irritum facient pactum meum 31:21 postquam inue-

bles are befallen them, that this song shall testify against them as a witness; for it shall not be forgotten out of the mouths of their seed: for I know their imagination which they go about, even now, before I have brought them into the land which I sware.

nerint eum mala multa et adflictiones respondebit ei canticum istud pro testimonio quod nulla delebit obliuio ex ore seminis tui scio enim cogitationes eius quae facturus sit hodie antequam introducam eum in terram quam ei pollicitus

31:22 Moses therefore wrote this song the same day, and sum 31:22 scripsit ergo Moses canticum et docuit filios Is-

taught it the children of Israel.

31:23 And he gave Joshua the son of Nun a charge, and rahel 31:23 praecepitque Iosue filio Nun et ait confortare

said, Be strong and of a good courage: for thou shalt bring the children of Israel into the land which I sware unto

them: and I will be with thee. et esto robustus tu enim introduces filios Israhel in terram 31:24 And it came to pass, when Moses had made an end quam pollicitus sum et ego ero tecum 31:24 postquam ergo

of writing the words of this law in a book, until they were finished,

scripsit Moses uerba legis huius in uolumine atque conpleuit

31:25 That Moses commanded the Levites, which bare the 31:25 praecepit Leuitis qui portabant arcam foederis Domini

ark of the covenant of the LORD, saying,

31:26 Take this book of the law, and put it in the side of dicens 31:26 tollite librum istum et ponite eum in latere ar-

the ark of the covenant of the LORD your God, that it may be there for a witness against thee.

cae foederis Domini Dei uestri ut sit ibi contra te in testi-

Liber Deuteronomii

monio 31:27 ego enim scio contentionem tuam et ceruicem 31:27 For I know thy rebellion, and thy stiff neck: behold, while I am yet alive with you this day, ye have been rebellious against the LORD; and how much more after my

tuam durissimam adhuc uiuente me et ingrediente uobiscum death? semper contentiose egistis contra Dominum quanto magis cum mortuus fuero 31:28 congregate ad me omnes maiores 31:28 Gather unto me all the elders of your tribes, and

your officers, that I may speak these words in their ears, and call heaven and earth to record against them.

natu per tribus uestras atque doctores et loquar audientibus eis sermones istos et inuocabo contra eos caelum et terram 31:29 noui enim quod post mortem meam inique agetis et de- 31:29 For I know that after my death ye will utterly cor-

rupt yourselves, and turn aside from the way which I have commanded you; and evil will befall you in the latter days; because ye will do evil in the sight of the LORD, to provoke him to anger through the work of your hands.

clinabitis cito de uia quam praecepi uobis et occurrent uobis mala in extremo tempore quando feceritis malum in conspectu Domini ut inritetis eum per opera manuum uestrarum 31:30 locutus est ergo Moses audiente uniuerso coetu Israhel 31:30 And Moses spake in the ears of all the congregation

of Israel the words of this song, until they were ended.

uerba carminis huius et ad finem usque conpleuit

32:1 audite caeli quae loquor audiat terra uerba oris mei 32:1 Give ear, O ye heavens, and I will speak; and hear, O

earth, the words of my mouth.

32:2 concrescat in pluuia doctrina mea fluat ut ros eloquium 32:2 My doctrine shall drop as the rain, my speech shall

distil as the dew, as the small rain upon the tender herb, and as the showers upon the grass:

meum quasi imber super herbam et quasi stillae super gra-

mina 32:3 quia nomen Domini inuocabo date magnificen- 32:3 Because I will publish the name of the LORD: ascribe

ye greatness unto our God.

tiam Deo nostro 32:4 Dei perfecta sunt opera et omnes uiae 32:4 He is the Rock, his work is perfect: for all his ways

are judgment: a God of truth and without iniquity, just and right is he.

eius iudicia Deus fidelis et absque ulla iniquitate iustus et rectus 32:5 peccauerunt ei non filii eius in sordibus gene- 32:5 They have corrupted themselves, their spot is not the

spot of his children: they are a perverse and crooked generation.

ratio praua atque peruersa 32:6 haecine reddis Domino po-

32:6 Do ye thus requite the LORD, O foolish people and unwise? is not he thy father that hath bought thee? hath he not made thee, and established thee?

pule stulte et insipiens numquid non ipse est pater tuus qui possedit et fecit et creauit te 32:7 memento dierum anti- 32:7 Remember the days of old, consider the years of

many generations: ask thy father, and he will shew thee; thy elders, and they will tell thee.

quorum cogita generationes singulas interroga patrem tuum et adnuntiabit tibi maiores tuos et dicent tibi 32:8 quando 32:8 When the Most High divided to the nations their in-

heritance, when he separated the sons of Adam, he set the bounds of the people according to the number of the chil-

diuidebat Altissimus gentes quando separabat filios Adam dren of Israel. constituit terminos populorum iuxta numerum filiorum Isra-

hel 32:9 pars autem Domini populus eius Iacob funiculus 32:9 For the LORD’s portion is his people; Jacob is the lot

of his inheritance.

hereditatis eius 32:10 inuenit eum in terra deserta in loco 32:10 He found him in a desert land, and in the waste

howling wilderness; he led him about, he instructed him, he kept him as the apple of his eye.

horroris et uastae solitudinis circumduxit eum et docuit et custodiuit quasi pupillam oculi sui 32:11 sicut aquila pro- 32:11 As an eagle stirreth up her nest, fluttereth over her

young, spreadeth abroad her wings, taketh them, beareth them on her wings:

uocans ad uolandum pullos suos et super eos uolitans expandit alas suas et adsumpsit eum atque portauit in umeris suis 32:12 Dominus solus dux eius fuit et non erat cum eo 32:12 So the LORD alone did lead him, and there was no

strange god with him.

Biblia Sacra Vulgata

32:13 He made him ride on the high places of the earth, deus alienus 32:13 constituit eum super excelsam terram ut

that he might eat the increase of the fields; and he made him to suck honey out of the rock, and oil out of the flinty

rock; comederet fructus agrorum ut sugeret mel de petra oleum-

32:14 Butter of kine, and milk of sheep, with fat of lambs, que de saxo durissimo 32:14 butyrum de armento et lac de and rams of the breed of Bashan, and goats, with the fat of kidneys of wheat; and thou didst drink the pure blood of

the grape. ouibus cum adipe agnorum et arietum filiorum Basan et hir-

cos cum medulla tritici et sanguinem uuae biberet meracissi-

32:15 But Jeshurun waxed fat, and kicked: thou art waxen mum 32:15 incrassatus est dilectus et recalcitrauit incrassa-

fat, thou art grown thick, thou art covered with fatness; then he forsook God which made him, and lightly es-

teemed the Rock of his salvation. tus inpinguatus dilatatus dereliquit Deum factorem suum et 32:16 They provoked him to jealousy with strange gods, recessit a Deo salutari suo 32:16 prouocauerunt eum in diis

with abominations provoked they him to anger.

alienis et in abominationibus ad iracundiam concitauerunt

32:17 They sacrificed unto devils, not to God; to gods 32:17 immolauerunt daemonibus et non Deo diis quos igno-

whom they knew not, to new gods that came newly up, whom your fathers feared not.

rabant noui recentesque uenerunt quos non coluerunt pat-

32:18 Of the Rock that begat thee thou art unmindful, and res eorum 32:18 Deum qui te genuit dereliquisti et oblitus

hast forgotten God that formed thee.

32:19 And when the LORD saw it, he abhorred them, be- es Domini creatoris tui 32:19 uidit Dominus et ad iracun-

cause of the provoking of his sons, and of his daughters.

diam concitatus est quia prouocauerunt eum filii sui et filiae

32:20 And he said, I will hide my face from them, I will 32:20 et ait abscondam faciem meam ab eis et considerabo

see what their end shall be: for they are a very froward generation, children in whom is no faith.

nouissima eorum generatio enim peruersa est et infideles filii

32:21 They have moved me to jealousy with that which is 32:21 ipsi me prouocauerunt in eo qui non erat Deus et inri-

not God; they have provoked me to anger with their vanities: and I will move them to jealousy with those which are not a people; I will provoke them to anger with a foolish nation.

tauerunt in uanitatibus suis et ego prouocabo eos in eo qui

32:22 For a fire is kindled in mine anger, and shall burn non est populus et in gente stulta inritabo illos 32:22 ignis

unto the lowest hell, and shall consume the earth with her increase, and set on fire the foundations of the mountains.

succensus est in furore meo et ardebit usque ad inferni nouissima deuorabitque terram cum germine suo et montium

32:23 I will heap mischiefs upon them; I will spend mine fundamenta conburet 32:23 congregabo super eos mala et

arrows upon them. 32:24 They shall be burnt with hunger, and devoured with sagittas meas conplebo in eis 32:24 consumentur fame et burning heat, and with bitter destruction: I will also send the teeth of beasts upon them, with the poison of serpents

of the dust. deuorabunt eos aues morsu amarissimo dentes bestiarum in-

mittam in eos cum furore trahentium super terram atque ser-

32:25 The sword without, and terror within, shall destroy pentium 32:25 foris uastabit eos gladius et intus pauor iuue-

both the young man and the virgin, the suckling also with the man of gray hairs.

nem simul ac uirginem lactantem cum homine sene 32:26

32:26 I said, I would scatter them into corners, I would make the remembrance of them to cease from among men:

dixi ubinam sunt cessare faciam ex hominibus memoriam

32:27 Were it not that I feared the wrath of the enemy, lest eorum 32:27 sed propter iram inimicorum distuli ne forte their adversaries should behave themselves strangely, and lest they should say, Our hand is high, and the LORD hath

not done all this. superbirent hostes eorum et dicerent manus nostra excelsa

32:28 For they are a nation void of counsel, neither is there et non Dominus fecit haec omnia 32:28 gens absque consiany understanding in them.

32:29 O that they were wise, that they understood this, that lio est et sine prudentia 32:29 utinam saperent et intellege-

they would consider their latter end!

32:30 How should one chase a thousand, and two put ten rent ac nouissima prouiderent 32:30 quomodo persequatur

thousand to flight, except their Rock had sold them, and the LORD had shut them up?

Liber Deuteronomii

unus mille et duo fugent decem milia nonne ideo quia Deus suus uendidit eos et Dominus conclusit illos 32:31 non 32:31 For their rock is not as our Rock, even our enemies

themselves being judges.

enim est Deus noster ut deus eorum et inimici nostri sunt iudices 32:32 de uinea Sodomorum uinea eorum et de su- 32:32 For their vine is of the vine of Sodom, and of the

fields of Gomorrah: their grapes are grapes of gall, their clusters are bitter:

burbanis Gomorrae uua eorum uua fellis et botri amarissimi 32:33 fel draconum uinum eorum et uenenum aspidum insa- 32:33 Their wine is the poison of dragons, and the cruel

venom of asps.

nabile 32:34 nonne haec condita sunt apud me et signata in 32:34 Is not this laid up in store with me, and sealed up

among my treasures?

thesauris meis 32:35 mea est ultio et ego retribuam in tem- 32:35 To me belongeth vengeance and recompence; their

foot shall slide in due time: for the day of their calamity is at hand, and the things that shall come upon them make

pore ut labatur pes eorum iuxta est dies perditionis et adesse haste. festinant tempora 32:36 iudicabit Dominus populum suum 32:36 For the LORD shall judge his people, and repent

himself for his servants, when he seeth that their power is gone, and there is none shut up, or left.

et in seruis suis miserebitur uidebit quod infirmata sit manus et clausi quoque defecerint residuique consumpti sint 32:37 et dicet ubi sunt dii eorum in quibus habebant fiduciam 32:37 And he shall say, Where are their gods, their rock in

whom they trusted,

32:38 de quorum uictimis comedebant adipes et bibebant ui- 32:38 Which did eat the fat of their sacrifices, and drank

the wine of their drink offerings? let them rise up and help you, and be your protection.

num libaminum surgant et opitulentur uobis et in necessitate uos protegant 32:39 uidete quod ego sim solus et non 32:39 See now that I, even I, am he, and there is no god

with me: I kill, and I make alive; I wound, and I heal: neither is there any that can deliver out of my hand.

sit alius deus praeter me ego occidam et ego uiuere faciam percutiam et ego sanabo et non est qui de manu mea possit eruere 32:40 leuabo ad caelum manum meam et dicam uiuo 32:40 For I lift up my hand to heaven, and say, I live for

ever.

ego in aeternum 32:41 si acuero ut fulgur gladium meum 32:41 If I whet my glittering sword, and mine hand take

hold on judgment; I will render vengeance to mine enemies, and will reward them that hate me.

et arripuerit iudicium manus mea reddam ultionem hostibus meis et his qui oderunt me retribuam 32:42 inebriabo sa- 32:42 I will make mine arrows drunk with blood, and my

sword shall devour flesh; and that with the blood of the slain and of the captives, from the beginning of revenges

gittas meas sanguine et gladius meus deuorabit carnes de upon the enemy. cruore occisorum et de captiuitate nudati inimicorum capitis 32:43 laudate gentes populum eius quia sanguinem seruorum 32:43 Rejoice, O ye nations, with his people: for he

will avenge the blood of his servants, and will render vengeance to his adversaries, and will be merciful unto

suorum ulciscetur et uindictam retribuet in hostes eorum et his land, and to his people. propitius erit terrae populi sui 32:44 uenit ergo Moses et lo- 32:44 And Moses came and spake all the words of this

song in the ears of the people, he, and Hoshea the son of Nun.

cutus est omnia uerba cantici huius in auribus populi ipse et Iosue filius Nun 32:45 conpleuitque omnes sermones istos 32:45 And Moses made an end of speaking all these words

to all Israel:

loquens ad uniuersum Israhel 32:46 et dixit ad eos ponite 32:46 And he said unto them, Set your hearts unto all the

words which I testify among you this day, which ye shall command your children to observe to do, all the words of

corda uestra in omnia uerba quae ego testificor uobis hodie this law. ut mandetis ea filiis uestris custodire et facere et implere uniuersa quae scripta sunt legis huius 32:47 quia non in cassum 32:47 For it is not a vain thing for you; because it is your

life: and through this thing ye shall prolong your days in the land, whither ye go over Jordan to possess it.

Biblia Sacra Vulgata

praecepta sunt uobis sed ut singuli in eis uiuerent quae facientes longo perseueretis tempore in terra ad quam Iordane

32:48 And the LORD spake unto Moses that selfsame day, transmisso ingredimini possidendam 32:48 locutusque est

saying, 32:49 Get thee up into this mountain Abarim, unto mount Dominus ad Mosen in eadem die dicens 32:49 ascende in Nebo, which is in the land of Moab, that is over against Jericho; and behold the land of Canaan, which I give unto

the children of Israel for a possession: montem istum Abarim id est transituum in montem Nebo qui est in terra Moab contra Hiericho et uide terram Chanaan quam ego tradam filiis Israhel obtinendam et morere in

32:50 And die in the mount whither thou goest up, and monte 32:50 quem conscendens iungeris populis tuis sicut

be gathered unto thy people; as Aaron thy brother died in mount Hor, and was gathered unto his people:

mortuus est Aaron frater tuus in monte Hor et adpositus po-

32:51 Because ye trespassed against me among the chil- pulis suis 32:51 quia praeuaricati estis contra me in medio

dren of Israel at the waters of MeribahKadesh, in the wilderness of Zin; because ye sanctified me not in the

midst of the children of Israel. filiorum Israhel ad aquas Contradictionis in Cades deserti

32:52 Yet thou shalt see the land before thee; but thou shalt Sin et non sanctificastis me inter filios Israhel 32:52 e connot go thither unto the land which I give the children of Israel.

tra uidebis terram et non ingredieris in eam quam ego dabo filiis Israhel

33:1 And this is the blessing, wherewith Moses the man of 33:1 haec est benedictio qua benedixit Moses homo Dei God blessed the children of Israel before his death.

33:2 And he said, The LORD came from Sinai, and rose filiis Israhel ante mortem suam 33:2 et ait Dominus de Sina

up from Seir unto them; he shined forth from mount Paran, and he came with ten thousands of saints: from his right

hand went a fiery law for them. uenit et de Seir ortus est nobis apparuit de monte Pharan et

33:3 Yea, he loved the people; all his saints are in thy hand: cum eo sanctorum milia in dextera eius ignea lex 33:3 diand they sat down at thy feet; every one shall receive of thy words.

lexit populos omnes sancti in manu illius sunt et qui adpro-

33:4 Moses commanded us a law, even the inheritance of pinquant pedibus eius accipient de doctrina illius 33:4 le-

the congregation of Jacob.

gem praecepit nobis Moses hereditatem multitudinis Iacob

33:5 And he was king in Jeshurun, when the heads of the 33:5 erit apud rectissimum rex congregatis principibus po-

people and the tribes of Israel were gathered together.

33:6 Let Reuben live, and not die; and let not his men be puli cum tribubus Israhel 33:6 uiuat Ruben et non moriatur

few.

33:7 And this is the blessing of Judah: and he said, Hear, et sit paruus in numero 33:7 haec est Iudae benedictio audi

LORD, the voice of Judah, and bring him unto his people: let his hands be sufficient for him; and be thou an help to

him from his enemies. Domine uocem Iudae et ad populum suum introduc eum ma-

nus eius pugnabunt pro eo et adiutor illius contra aduersarios

33:8 And of Levi he said, Let thy Thummim and thy Urim eius erit 33:8 Leui quoque ait perfectio tua et doctrina tua

be with thy holy one, whom thou didst prove at Massah, and with whom thou didst strive at the waters of Meribah;

uiro sancto tuo quem probasti in Temptatione et iudicasti

33:9 Who said unto his father and to his mother, I have ad aquas Contradictionis 33:9 qui dixit patri suo et matri

not seen him; neither did he acknowledge his brethren, nor knew his own children: for they have observed thy word,

and kept thy covenant. suae nescio uos et fratribus suis ignoro illos et nescierunt

filios suos hii custodierunt eloquium tuum et pactum tuum

33:10 They shall teach Jacob thy judgments, and Israel thy seruauerunt 33:10 iudicia tua o Iacob et legem tuam o Islaw: they shall put incense before thee, and whole burnt sacrifice upon thine altar.

rahel ponent thymiama in furore tuo et holocaustum super

Liber Deuteronomii

altare tuum 33:11 benedic Domine fortitudini eius et opera 33:11 Bless, LORD, his substance, and accept the work of

his hands; smite through the loins of them that rise against him, and of them that hate him, that they rise not again.

manuum illius suscipe percute dorsa inimicorum eius et qui oderunt eum non consurgant 33:12 et Beniamin ait aman- 33:12 And of Benjamin he said, The beloved of the LORD

shall dwell in safety by him; and the Lord shall cover him all the day long, and he shall dwell between his shoulders.

tissimus Domini habitabit confidenter in eo quasi in thalamo tota die morabitur et inter umeros illius requiescet 33:13 Io- 33:13 And of Joseph he said, Blessed of the LORD be his

land, for the precious things of heaven, for the dew, and for the deep that coucheth beneath,

seph quoque ait de benedictione Domini terra eius de po-

mis caeli et rore atque abysso subiacente 33:14 de pomis 33:14 And for the precious fruits brought forth by the sun,

and for the precious things put forth by the moon,

fructuum solis ac lunae 33:15 de uertice antiquorum mon- 33:15 And for the chief things of the ancient mountains,

and for the precious things of the lasting hills,

tium de pomis collium aeternorum 33:16 et de frugibus ter- 33:16 And for the precious things of the earth and fulness

thereof, and for the good will of him that dwelt in the bush: let the blessing come upon the head of Joseph, and upon the top of the head of him that was separated from his brethren.

rae et plenitudine eius benedictio illius qui apparuit in rubo ueniat super caput Ioseph et super uerticem nazarei inter fratres suos 33:17 quasi primogeniti tauri pulchritudo eius cor- 33:17 His glory is like the firstling of his bullock, and his

horns are like the horns of unicorns: with them he shall push the people together to the ends of the earth: and they are the ten thousands of Ephraim, and they are the thousands of Manasseh.

nua rinocerotis cornua illius in ipsis uentilabit gentes usque ad terminos terrae hae sunt multitudines Ephraim et haec milia Manasse 33:18 et Zabulon ait laetare Zabulon in exitu 33:18 And of Zebulun he said, Rejoice, Zebulun, in thy

going out; and, Issachar, in thy tents.

tuo et Isachar in tabernaculis tuis 33:19 populos ad mon- 33:19 They shall call the people unto the mountain; there

they shall offer sacrifices of righteousness: for they shall suck of the abundance of the seas, and of treasures hid in

tem uocabunt ibi immolabunt uictimas iustitiae qui inunda- the sand. tionem maris quasi lac sugent et thesauros absconditos harenarum 33:20 et Gad ait benedictus in latitudine Gad quasi 33:20 And of Gad he said, Blessed be he that enlargeth

Gad: he dwelleth as a lion, and teareth the arm with the crown of the head.

leo requieuit cepitque brachium et uerticem 33:21 et uidit

33:21 And he provided the first part for himself, because there, in a portion of the lawgiver, was he seated; and he came with the heads of the people, he executed the justice of the LORD, and his judgments with Israel.

principatum suum quod in parte sua doctor esset repositus qui fuit cum principibus populi et fecit iustitias Domini et iudicium suum cum Israhel 33:22 Dan quoque ait Dan ca- 33:22 And of Dan he said, Dan is a lion’s whelp: he shall

leap from Bashan.

tulus leonis fluet largiter de Basan 33:23 et Nepthalim dixit 33:23 And of Naphtali he said, O Naphtali, satisfied with

favour, and full with the blessing of the LORD: possess thou the west and the south.

Nepthalim abundantia perfruetur et plenus erit benedictione Domini mare et meridiem possidebit 33:24 Aser quoque ait 33:24 And of Asher he said, Let Asher be blessed with

children; let him be acceptable to his brethren, and let him dip his foot in oil.

benedictus in filiis Aser sit placens fratribus suis tinguat in

oleo pedem suum 33:25 ferrum et aes calciamentum eius 33:25 Thy shoes shall be iron and brass; and as thy days,

so shall thy strength be.

sicut dies iuuentutis tuae ita et senectus tua 33:26 non est 33:26 There is none like unto the God of Jeshurun, who

rideth upon the heaven in thy help, and in his excellency on the sky.

alius ut Deus rectissimi ascensor caeli auxiliator tuus magnificentia eius discurrunt nubes 33:27 habitaculum eius sur- 33:27 The eternal God is thy refuge, and underneath are

the everlasting arms: and he shall thrust out the enemy from before thee; and shall say, Destroy them.

sum et subter brachia sempiterna eiciet a facie tua inimicum dicetque conterere 33:28 habitabit Israhel confidenter et so- 33:28 Israel then shall dwell in safety alone: the fountain

of Jacob shall be upon a land of corn and wine; also his heavens shall drop down dew.

Biblia Sacra Vulgata

lus oculus Iacob in terra frumenti et uini caelique caligabunt

33:29 Happy art thou, O Israel: who is like unto thee, O rore 33:29 beatus tu Israhel quis similis tui popule qui sal-

people saved by the LORD, the shield of thy help, and who is the sword of thy excellency! and thine enemies shall be found liars unto thee; and thou shalt tread upon their high places.

uaris in Domino scutum auxilii tui et gladius gloriae tuae negabunt te inimici tui et tu eorum colla calcabis

34:1 And Moses went up from the plains of Moab unto 34:1 ascendit ergo Moses de campestribus Moab super

the mountain of Nebo, to the top of Pisgah, that is over against Jericho. And the LORD shewed him all the land

of Gilead, unto Dan, montem Nebo in uerticem Phasga contra Hiericho ostendit-

34:2 And all Naphtali, and the land of Ephraim, and Man- que ei Dominus omnem terram Galaad usque Dan 34:2 et asseh, and all the land of Judah, unto the utmost sea,

uniuersum Nepthalim terramque Ephraim et Manasse et om-

34:3 And the south, and the plain of the valley of Jericho, nem terram usque ad mare Nouissimum 34:3 et australem the city of palm trees, unto Zoar.

partem et latitudinem campi Hiericho ciuitatis Palmarum us-

34:4 And the LORD said unto him, This is the land which que Segor 34:4 dixitque Dominus ad eum haec est terra pro

I sware unto Abraham, unto Isaac, and unto Jacob, saying, I will give it unto thy seed: I have caused thee to see it

with thine eyes, but thou shalt not go over thither. qua iuraui Abraham Isaac et Iacob dicens semini tuo dabo 34:5 So Moses the servant of the LORD died there in the eam uidisti eam oculis tuis et non transibis ad illam 34:5

land of Moab, according to the word of the LORD.

mortuusque est ibi Moses seruus Domini in terra Moab iu-

34:6 And he buried him in a valley in the land of Moab, bente Domino 34:6 et sepeliuit eum in ualle terrae Moab

over against Bethpeor: but no man knoweth of his sepulchre unto this day.

contra Phogor et non cognouit homo sepulchrum eius usque

34:7 And Moses was an hundred and twenty years old in praesentem diem 34:7 Moses centum et uiginti annorum

when he died: his eye was not dim, nor his natural force abated.

erat quando mortuus est non caligauit oculus eius nec dentes

34:8 And the children of Israel wept for Moses in the illius moti sunt 34:8 fleueruntque eum filii Israhel in cam-

plains of Moab thirty days: so the days of weeping and mourning for Moses were ended.

pestribus Moab triginta diebus et conpleti sunt dies planctus

34:9 And Joshua the son of Nun was full of the spirit lugentium Mosen 34:9 Iosue uero filius Nun repletus est

of wisdom; for Moses had laid his hands upon him: and the children of Israel hearkened unto him, and did as the

LORD commanded Moses. spiritu sapientiae quia Moses posuit super eum manus suas

et oboedierunt ei filii Israhel feceruntque sicut praecepit Do-

34:10 And there arose not a prophet since in Israel like minus Mosi 34:10 et non surrexit propheta ultra in Israhel

unto Moses, whom the LORD knew face to face,

34:11 In all the signs and the wonders, which the LORD sicut Moses quem nosset Dominus facie ad faciem 34:11 in

sent him to do in the land of Egypt to Pharaoh, and to all his servants, and to all his land,

omnibus signis atque portentis quae misit per eum ut faceret in terra Aegypti Pharaoni et omnibus seruis eius uniuersae-

34:12 And in all that mighty hand, and in all the great que terrae illius 34:12 et cunctam manum robustam magna-

terror which Moses shewed in the sight of all Israel.

\underline{Liber Iosue}

que mirabilia quae fecit Moses coram uniuerso Israhel 1:1 et factum est ut post mortem Mosi serui Domini loque- 1:1 Now after the death of Moses the servant of the LORD

it came to pass, that the LORD spake unto Joshua the son of Nun, Moses’ minister, saying,

retur Dominus ad Iosue filium Nun ministrum Mosi et dice-

ret ei 1:2 Moses seruus meus mortuus est surge et transi 1:2 Moses my servant is dead; now therefore arise, go over this Jordan, thou, and all this people, unto the land which I do give to them, even to the children of Israel.

Iordanem istum tu et omnis populus tecum in terram quam

ego dabo filiis Israhel 1:3 omnem locum quem calcauerit 1:3 Every place that the sole of your foot shall tread upon,

that have I given unto you, as I said unto Moses.

uestigium pedis uestri uobis tradam sicut locutus sum Mosi

1:4 a deserto et Libano usque ad fluuium magnum Eufra- 1:4 From the wilderness and this Lebanon even unto the great river, the river Euphrates, all the land of the Hittites, and unto the great sea toward the going down of the sun,

ten omnis terra Hettheorum usque ad mare Magnum contra shall be your coast. solis occasum erit terminus uester 1:5 nullus uobis poterit 1:5 There shall not any man be able to stand before thee

all the days of thy life: as I was with Moses, so I will be with thee: I will not fail thee, nor forsake thee.

resistere cunctis diebus uitae tuae sicut fui cum Mose ero et tecum non dimittam nec derelinquam te 1:6 confortare et 1:6 Be strong and of a good courage: for unto this people

shalt thou divide for an inheritance the land, which I sware unto their fathers to give them.

esto robustus tu enim sorte diuides populo huic terram pro qua iuraui patribus suis ut traderem eam illis 1:7 confortare 1:7 Only be thou strong and very courageous, that thou

mayest observe to do according to all the law, which Moses my servant commanded thee: turn not from it to the right hand or to the left, that thou mayest prosper withersoever thou goest.

igitur et esto robustus ualde ut custodias et facias omnem legem quam praecepit tibi Moses seruus meus ne declines ab ea ad dextram uel ad sinistram ut intellegas cuncta quae agis 1:8 non recedat uolumen legis huius de ore tuo sed medita- 1:8 This book of the law shall not depart out of thy

mouth; but thou shalt meditate therein day and night, that thou mayest observe to do according to all that is written therein: for then thou shalt make thy way prosperous, and then thou shalt have good success.

beris in eo diebus ac noctibus ut custodias et facias omnia quae scripta sunt in eo tunc diriges uiam tuam et intelleges eam 1:9 ecce praecipio tibi confortare et esto robustus noli 1:9 Have not I commanded thee? Be strong and of a good

courage; be not afraid, neither be thou dismayed: for the LORD thy God is with thee whithersoever thou goest.

metuere et noli timere quoniam tecum est Dominus Deus tuus in omnibus ad quaecumque perrexeris 1:10 praecepit- 1:10 Then Joshua commanded the officers of the people,

saying,

que Iosue principibus populi dicens transite per medium castrorum et imperate populo ac dicite 1:11 praeparate uobis 1:11 Pass through the host, and command the people, say-

ing, Prepare you victuals; for within three days ye shall pass over this Jordan, to go in to possess the land, which

cibaria quoniam post diem tertium transibitis Iordanem et in- the LORD your God giveth you to possess it. trabitis ad possidendam terram quam Dominus Deus uester daturus est uobis 1:12 Rubenitis quoque et Gadditis et di- 1:12 And to the Reubenites, and to the Gadites, and to half

the tribe of Manasseh, spake Joshua, saying,

midiae tribui Manasse ait 1:13 mementote sermonis quem 1:13 Remember the word which Moses the servant of the

LORD commanded you, saying, The LORD your God hath given you rest, and hath given you this land.

praecepit uobis Moses famulus Domini dicens Dominus Deus uester dedit uobis requiem et omnem terram 1:14 uxores 1:14 Your wives, your little ones, and your cattle, shall

remain in the land which Moses gave you on this side Jordan; but ye shall pass before your brethren armed, all the

uestrae et filii ac iumenta manebunt in terra quam tradidit mighty men of valour, and help them; uobis Moses trans Iordanem uos autem transite armati ante fratres uestros omnes fortes manu et pugnate pro eis 1:15 1:15 Until the LORD have given your brethren rest, as

he hath given you, and they also have possessed the land which the LORD your God giveth them: then ye shall return unto the land of your possession, and enjoy it, which Moses the LORD’s servant gave you on this side Jordan toward the sunrising.

Biblia Sacra Vulgata

donec det requiem Dominus fratribus uestris sicut et uobis dedit et possideant ipsi quoque terram quam Dominus Deus uester daturus est eis et sic reuertemini in terram possessionis uestrae et habitabitis in ea quam uobis dedit Moses famu-

1:16 And they answered Joshua, saying, All that thou lus Domini trans Iordanem contra solis ortum 1:16 respon-

commandest us we will do, and whithersoever thou sendest us, we will go.

deruntque ad Iosue atque dixerunt omnia quae praecepisti

1:17 According as we hearkened unto Moses in all things, nobis faciemus et quocumque miseris ibimus 1:17 sicut ob-

so will we hearken unto thee: only the LORD thy God be with thee, as he was with Moses.

oediuimus in cunctis Mosi ita oboediemus et tibi tantum sit

1:18 Whosoever he be that doth rebel against thy com- Dominus Deus tecum sicut fuit cum Mose 1:18 qui conmandment, and will not hearken unto thy words in all that thou commandest him, he shall be put to death: only be

strong and of a good courage. tradixerit ori tuo et non oboedierit cunctis sermonibus quos praeceperis ei moriatur tu tantum confortare et uiriliter age

2:1 And Joshua the son of Nun sent out of Shittim two men 2:1 misit ergo Iosue filius Nun de Setthim duos uiros ex-

to spy secretly, saying, Go view the land, even Jericho. And they went, and came into an harlot’s house, named

Rahab, and lodged there. ploratores abscondito et dixit eis ite et considerate terram

urbemque Hiericho qui pergentes ingressi sunt domum mu-

2:2 And it was told the king of Jericho, saying, Behold, lieris meretricis nomine Raab et quieuerunt apud eam 2:2

there came men in hither to night of the children of Israel to search out the country.

nuntiatumque est regi Hiericho et dictum ecce uiri ingressi sunt huc per noctem de filiis Israhel ut explorarent terram

2:3 And the king of Jericho sent unto Rahab, saying, Bring 2:3 misitque rex Hiericho ad Raab dicens educ uiros qui forth the men that are come to thee, which are entered into thine house: for they be come to search out all the country.

uenerunt ad te et ingressi sunt domum tuam exploratores

2:4 And the woman took the two men, and hid them, and quippe sunt et omnem terram considerare uenerunt 2:4 tol-

said thus, There came men unto me, but I wist not whence they were:

lensque mulier uiros abscondit et ait fateor uenerunt ad me

2:5 And it came to pass about the time of shutting of the sed nesciebam unde essent 2:5 cumque porta clauderetur

gate, when it was dark, that the men went out: whither the men went I wot not: pursue after them quickly; for ye

shall overtake them. in tenebris et illi pariter exierunt nescio quo abierunt per-2:6 But she had brought them up to the roof of the house, sequimini cito et conprehendetis eos 2:6 ipsa autem fecit

and hid them with the stalks of flax, which she had laid in order upon the roof.

ascendere uiros in solarium domus suae operuitque eos lini

2:7 And the men pursued after them the way to Jordan stipula quae ibi erat 2:7 hii autem qui missi fuerant secuti

unto the fords: and as soon as they which pursued after them were gone out, they shut the gate.

sunt eos per uiam quae ducit ad uadum Iordanis illisque eg-

2:8 And before they were laid down, she came up unto ressis statim porta clausa est 2:8 necdum obdormierant qui

them upon the roof;

2:9 And she said unto the men, I know that the LORD hath latebant et ecce mulier ascendit ad eos et ait 2:9 noui quod

given you the land, and that your terror is fallen upon us, and that all the inhabitants of the land faint because of you.

tradiderit Dominus uobis terram etenim inruit in nos terror

2:10 For we have heard how the LORD dried up the water uester et elanguerunt omnes habitatores terrae 2:10 audiui-

of the Red sea for you, when ye came out of Egypt; and what ye did unto the two kings of the Amorites, that were on the other side Jordan, Sihon and Og, whom ye utterly destroyed.

mus quod siccauerit Dominus aquas maris Rubri ad uestrum introitum quando egressi estis ex Aegypto et quae feceritis duobus Amorreorum regibus qui erant trans Iordanem Seon et Og quos interfecistis 2:11 et haec audientes pertimuimus 2:11 And as soon as we had heard these things, our hearts

did melt, neither did there remain any more courage in any man, because of you: for the LORD your God, he is God

et elanguit cor nostrum nec remansit in nobis spiritus ad in- in heaven above, and in earth beneath. troitum uestrum Dominus enim Deus uester ipse est Deus in caelo sursum et in terra deorsum 2:12 nunc ergo iurate 2:12 Now therefore, I pray you, swear unto me by the

LORD, since I have shewed you kindness, that ye will also shew kindness unto my father’s house, and give me a true

mihi per Dominum ut quomodo ego feci uobiscum mise- token: ricordiam ita et uos faciatis cum domo patris mei detisque

mihi signum uerum 2:13 et saluetis patrem meum et ma- 2:13 And that ye will save alive my father, and my mother, and my brethren, and my sisters, and all that they have, and deliver our lives from death.

trem fratres ac sorores meas et omnia quae eorum sunt et eruatis animas nostras de morte 2:14 qui responderunt ei 2:14 And the men answered her, Our life for yours, if

ye utter not this our business. And it shall be, when the LORD hath given us the land, that we will deal kindly and

anima nostra sit pro uobis in mortem si tamen non prodiieris truly with thee. nos cumque tradiderit nobis Dominus terram faciemus in te misericordiam et ueritatem 2:15 dimisit ergo eos per funem 2:15 Then she let them down by a cord through the win-

dow: for her house was upon the town wall, and she dwelt upon the wall.

de fenestra domus enim eius herebat muro 2:16 dixitque ad

2:16 And she said unto them, Get you to the mountain, lest the pursuers meet you; and hide yourselves there three days, until the pursuers be returned: and afterward may ye go your way.

eos ad montana conscendite ne forte occurrant uobis reuertentes ibique latete diebus tribus donec redeent et sic ibitis per uiam uestram 2:17 qui dixerunt ad eam innoxii erimus a 2:17 And the men said unto her, We will be blameless of

this thine oath which thou hast made us swear.

iuramento hoc quo adiurasti nos 2:18 si ingredientibus no- 2:18 Behold, when we come into the land, thou shalt bind

this line of scarlet thread in the window which thou didst let us down by: and thou shalt bring thy father, and thy mother, and thy brethren, and all thy father’s household, home unto thee.

bis terram signum fuerit funiculus iste coccineus et ligaueris eum in fenestra per quam nos dimisisti et patrem tuum ac matrem fratresque et omnem cognationem tuam congregaueris in domum tuam 2:19 qui ostium domus tuae egressus 2:19 And it shall be, that whosoever shall go out of the

doors of thy house into the street, his blood shall be upon his head, and we will be guiltless: and whosoever shall be with thee in the house, his blood shall be on our head, if any hand be upon him.

fuerit sanguis ipsius erit in caput eius et nos erimus alieni cunctorum autem sanguis qui tecum fuerint in domo redundabit in caput nostrum si eos aliquis tetigerit 2:20 quod si 2:20 And if thou utter this our business, then we will be

quit of thine oath which thou hast made us to swear.

nos prodeee uolueris et sermonem istum proferre in medium erimus mundi ab hoc iuramento quo adiurasti nos 2:21 et 2:21 And she said, According unto your words, so be it.

And she sent them away, and they departed: and she bound the scarlet line in the window.

illa respondit sicut locuti estis ita fiat dimittensque eos ut pergerent adpendit funiculum coccineum in fenestra 2:22 2:22 And they went, and came unto the mountain, and

abode there three days, until the pursuers were returned: and the pursuers sought them throughout all the way, but

illi uero ambulantes peruenerunt ad montana et manserunt found them not. ibi tres dies donec reuerterentur qui fuerant persecuti quaerentes enim per omnem uiam non reppererunt eos 2:23 qui- 2:23 So the two men returned, and descended from the

mountain, and passed over, and came to Joshua the son of Nun, and told him all things that befell them:

bus urbem ingressis reuersi sunt et descenderunt explorato-

Biblia Sacra Vulgata

res de monte et Iordane transmisso uenerunt ad Iosue filium

2:24 And they said unto Joshua, Truly the LORD hath de- Nun narraueruntque ei omnia quae acciderant sibi 2:24 at-

livered into our hands all the land; for even all the inhabitants of the country do faint because of us.

que dixerunt tradidit Dominus in manus nostras omnem terram hanc et timore prostrati sunt cuncti habitatores eius

3:1 And Joshua rose early in the morning; and they re- 3:1 igitur Iosue de nocte consurgens mouit castra egre-

moved from Shittim, and came to Jordan, he and all the children of Israel, and lodged there before they passed

over. dientesque de Setthim uenerunt ad Iordanem ipse et om-3:2 And it came to pass after three days, that the officers nes filii Israhel et morati sunt ibi per tres dies 3:2 quibus

went through the host;

3:3 And they commanded the people, saying, When ye see euolutis transierunt praecones per castrorum medium 3:3

the ark of the covenant of the LORD your God, and the priests the Levites bearing it, then ye shall remove from

your place, and go after it. et clamare coeperunt quando uideritis arcam foederis Do-

mini Dei uestri et sacerdotes stirpis leuiticae portantes eam

3:4 Yet there shall be a space between you and it, about uos quoque consurgite et sequimini praecedentes 3:4 sitque

two thousand cubits by measure: come not near unto it, that ye may know the way by which ye must go: for ye

have not passed this way heretofore. inter uos et arcam spatium cubitorum duum milium ut procul uidere possitis et nosse per quam uiam ingrediamini quia prius non ambulastis per eam et cauete ne adpropinquetis ad

3:5 And Joshua said unto the people, Sanctify yourselves: arcam 3:5 dixitque Iosue ad populum sanctificamini cras for to morrow the LORD will do wonders among you.

3:6 And Joshua spake unto the priests, saying, Take up the enim faciet Dominus inter uos mirabilia 3:6 et ait ad sacer-

ark of the covenant, and pass over before the people. And they took up the ark of the covenant, and went before the

people. dotes tollite arcam foederis et praecedite populum qui iussa 3:7 And the LORD said unto Joshua, This day will I begin conplentes tulerunt et ambulauerunt ante eos 3:7 dixitque

to magnify thee in the sight of all Israel, that they may know that, as I was with Moses, so I will be with thee.

Dominus ad Iosue hodie incipiam exaltare te coram omni Israhel ut sciant quod sicut cum Mosi fui ita et tecum sim

3:8 And thou shalt command the priests that bear the ark 3:8 tu autem praecipe sacerdotibus qui portant arcam foe-

of the covenant, saying, When ye are come to the brink of the water of Jordan, ye shall stand still in Jordan.

deris et dic eis cum ingressi fueritis partem aquae Iordanis

3:9 And Joshua said unto the children of Israel, Come state in ea 3:9 dixitque Iosue ad filios Israhel accedite huc

hither, and hear the words of the LORD your God. 3:10 And Joshua said, Hereby ye shall know that the living et audite uerba Domini Dei uestri 3:10 et rursum in hoc in-God is among you, and that he will without fail drive out from before you the Canaanites, and the Hittites, and the Hivites, and the Perizzites, and the Girgashites, and the Amorites, and the Jebusites.

quit scietis quod Dominus Deus uiuens in medio uestri est et disperdat in conspectu uestro Chananeum Hettheum Eueum et Ferezeum Gergeseum quoque et Amorreum et Iebuseum

3:11 Behold, the ark of the covenant of the LORD of all 3:11 ecce arca foederis Domini omnis terrae antecedet uos

the earth passeth over before you into Jordan.

3:12 Now therefore take you twelve men out of the tribes per Iordanem 3:12 parate duodecim uiros de tribubus Isra-

of Israel, out of every tribe a man.

3:13 And it shall come to pass, as soon as the soles of hel singulos per singulas tribus 3:13 et cum posuerint uesti-

the feet of the priests that bear the ark of the LORD, the LORD of all the earth, shall rest in the waters of Jordan, that the waters of Jordan shall be cut off from the waters that come down from above; and they shall stand upon an heap.

gia pedum suorum sacerdotes qui portant arcam Domini Dei uniuersae terrae in aquis Iordanis aquae quae inferiores sunt decurrent atque deficient quae autem desuper ueniunt in una mole consistent 3:14 igitur egressus est populus de taberna- 3:14 And it came to pass, when the people removed from

their tents, to pass over Jordan, and the priests bearing the ark of the covenant before the people;

culis suis ut transirent Iordanem et sacerdotes qui portabant arcam foederis pergebant ante eum 3:15 ingressisque eis 3:15 And as they that bare the ark were come unto Jordan,

and the feet of the priests that bare the ark were dipped in the brim of the water, (for Jordan overfloweth all his banks

Iordanem et pedibus eorum tinctis in parte aquae cum Ior- all the time of harvest,) danis autem ripas aluei sui tempore messis impleret 3:16 3:16 That the waters which came down from above stood

and rose up upon an heap very far from the city Adam, that is beside Zaretan: and those that came down toward the sea of the plain, even the salt sea, failed, and were cut off: and the people passed over right against Jericho.

steterunt aquae descendentes in uno loco et instar montis intumescentes apparebant procul ab urbe quae uocatur Adom usque ad locum Sarthan quae autem inferiores erant in mare Solitudinis quod nunc uocatur Mortuum descenderunt usquequo omnino deficerent 3:17 populus autem incedebat 3:17 And the priests that bare the ark of the covenant of

the LORD stood firm on dry ground in the midst of Jordan, and all the Israelites passed over on dry ground, until all

contra Iordanem et sacerdotes qui portabant arcam foederis the people were passed clean over Jordan. Domini stabant super siccam humum in medio Iordanis accincti omnisque populus per arentem alueum transiebat

4:1 quibus transgressis dixit Dominus ad Iosue 4:2 elige 4:1 And it came to pass, when all the people were clean

passed over Jordan, that the LORD spake unto Joshua, saying, 4:2 Take you twelve men out of the people, out of every tribe a man,

duodecim uiros singulos per singulas tribus 4:3 et praecipe

4:3 And command ye them, saying, Take you hence out of the midst of Jordan, out of the place where the priests’ feet stood firm, twelve stones, and ye shall carry them over with you, and leave them in the lodging place, where ye shall lodge this night.

eis ut tollant de medio Iordanis alueo ubi steterunt sacerdotum pedes duodecim durissimos lapides quos ponetis in loco castrorum ubi fixeritis hac nocte tentoria 4:4 uocauitque Io- 4:4 Then Joshua called the twelve men, whom he had pre-

pared of the children of Israel, out of every tribe a man:

sue duodecim uiros quos elegerat de filiis Israhel singulos de tribubus singulis 4:5 et ait ad eos ite ante arcam Domini Dei 4:5 And Joshua said unto them, Pass over before the ark of

the LORD your God into the midst of Jordan, and take you up every man of you a stone upon his shoulder, according

uestri ad Iordanis medium et portate singuli singulos lapides unto the number of the tribes of the children of Israel: in umeris uestris iuxta numerum filiorum Israhel 4:6 ut sit 4:6 That this may be a sign among you, that when your

children ask their fathers in time to come, saying, What mean ye by these stones?

signum inter uos et quando interrogauerint uos filii uestri cras dicentes quid sibi uolunt isti lapides 4:7 respondebi- 4:7 Then ye shall answer them, That the waters of Jordan

were cut off before the ark of the covenant of the LORD; when it passed over Jordan, the waters of Jordan were cut off: and these stones shall be for a memorial unto the children of Israel for ever.

tis eis defecerunt aquae Iordanis ante arcam foederis Domini cum transiret eum idcirco positi sunt lapides isti in monumentum filiorum Israhel usque in aeternum 4:8 fecerunt 4:8 And the children of Israel did so as Joshua com-

manded, and took up twelve stones out of the midst of Jordan, as the LORD spake unto Joshua, according to the number of the tribes of the children of Israel, and carried them over with them unto the place where they lodged, and laid them down there.

ergo filii Israhel sicut eis praecepit Iosue portantes de medio Iordanis alueo duodecim lapides ut ei Dominus imperarat iuxta numerum filiorum Israhel usque ad locum in quo castrametati sunt ibique posuerunt eos 4:9 alios quoque duo- 4:9 And Joshua set up twelve stones in the midst of Jordan,

in the place where the feet of the priests which bare the ark of the covenant stood: and they are there unto this day.

decim lapides posuit Iosue in medio Iordanis alueo ubi steterunt sacerdotes qui portabant arcam foederis et sunt ibi us-

Biblia Sacra Vulgata

4:10 For the priests which bare the ark stood in the midst que in praesentem diem 4:10 sacerdotes autem qui portaof Jordan, until everything was finished that the LORD commanded Joshua to speak unto the people, according to all that Moses commanded Joshua: and the people hasted and passed over.

bant arcam stabant in Iordanis medio donec omnia conplerentur quae Iosue ut loqueretur ad populum praeceperat Dominus et dixerat ei Moses festinauitque populus et transiit

4:11 And it came to pass, when all the people were clean 4:11 cumque transissent omnes transiuit et arca Domini sa-

passed over, that the ark of the LORD passed over, and the priests, in the presence of the people.

cerdotesque pergebant ante populum 4:12 filii quoque Ru-

4:12 And the children of Reuben, and the children of Gad, and half the tribe of Manasseh, passed over armed before the children of Israel, as Moses spake unto them:

ben et Gad et dimidiae tribus Manasse armati praecedebant

4:13 About forty thousand prepared for war passed over filios Israhel sicut eis praeceperat Moses 4:13 et quadra-

before the LORD unto battle, to the plains of Jericho.

ginta pugnatorum milia per turmas et cuneos incedebant per

4:14 On that day the LORD magnified Joshua in the sight plana atque campestria urbis Hiericho 4:14 in illo die mag-

of all Israel; and they feared him, as they feared Moses, all the days of his life.

nificauit Dominus Iosue coram omni Israhel ut timerent eum

4:15 And the LORD spake unto Joshua, saying, sicut timuerant Mosen dum aduiueret 4:15 dixitque ad eum 4:16 Command the priests that bear the ark of the testi- 4:16 praecipe sacerdotibus qui portant arcam foederis ut as-

mony, that they come up out of Jordan.

4:17 Joshua therefore commanded the priests, saying, cendant de Iordane 4:17 qui praecepit eis dicens ascendite

Come ye up out of Jordan.

4:18 And it came to pass, when the priests that bare the de Iordane 4:18 cumque ascendissent portantes arcam foe-

ark of the covenant of the LORD were come up out of the midst of Jordan, and the soles of the priests’ feet were lifted up unto the dry land, that the waters of Jordan returned unto their place, and flowed over all his banks, as they did before.

deris Domini et siccam humum calcare coepissent reuersae sunt aquae in alueum suum et fluebant sicut ante consueue-

4:19 And the people came up out of Jordan on the tenth rant 4:19 populus autem ascendit de Iordane decimo men-

day of the first month, and encamped in Gilgal, in the east border of Jericho.

sis primi die et castrametati sunt in Galgalis contra orienta-

4:20 And those twelve stones, which they took out of Jor- lem plagam urbis Hiericho 4:20 duodecim quoque lapides dan, did Joshua pitch in Gilgal.

quos de Iordanis alueo sumpserant posuit Iosue in Galga-

4:21 And he spake unto the children of Israel, saying, lis 4:21 et dixit ad filios Israhel quando interrogauerint filii

When your children shall ask their fathers in time to come, saying, What mean these stones?

uestri cras patres suos et dixerint eis quid sibi uolunt isti la-

4:22 Then ye shall let your children know, saying, Israel pides 4:22 docebitis eos atque dicetis per arentem alueum

came over this Jordan on dry land.

4:23 For the LORD your God dried up the waters of Jordan transiuit Israhel Iordanem istum 4:23 siccante Domino Deo

from before you, until ye were passed over,

4:24 as the LORD your God did to the Red sea, which he uestro aquas eius in conspectu uestro donec transiretis 4:24

dried up from before us, until we were gone over:

sicut fecerat prius in mari Rubro quod siccauit donec trans-

4:25 That all the people of the earth might know the hand iremus 4:25 ut discant omnes terrarum populi fortissimam

of the LORD, that it is mighty: that ye might fear the LORD your God for ever.

Domini manum et ut uos timeatis Dominum Deum uestrum omni tempore

5:1 And it came to pass, when all the kings of the Amor- 5:1 postquam ergo audierunt omnes reges Amorreorum ites, which were on the side of Jordan westward, and all the kings of the Canaanites, which were by the sea, heard that the LORD had dried up the waters of Jordan from before the children of Israel, until we were passed over, that their heart melted, neither was there spirit in them any more, because of the children of Israel.

qui habitabant trans Iordanem ad occidentalem plagam et cuncti reges Chanaan qui propinqua possidebant Magno mari loca quod siccasset Dominus fluenta Iordanis coram filiis Israhel donec transirent dissolutum est cor eorum et non remansit in eis spiritus timentium introitum filiorum Israhel 5:2 eo tempore ait Dominus ad Iosue fac tibi cultros lapideos 5:2 At that time the LORD said unto Joshua, Make thee

sharp knives, and circumcise again the children of Israel the second time.

et circumcide secundo filios Israhel 5:3 fecit quod iusse-

5:3 And Joshua made him sharp knives, and circumcised the children of Israel at the hill of the foreskins.

rat Dominus et circumcidit filios Israhel in colle Praeputiorum 5:4 haec autem causa est secundae circumcisionis om- 5:4 And this is the cause why Joshua did circumcise: All

the people that came out of Egypt, that were males, even all the men of war, died in the wilderness by the way, after

nis populus qui egressus est ex Aegypto generis masculini they came out of Egypt. uniuersi bellatores uiri mortui sunt in deserto per longissimos uiae circuitus 5:5 qui omnes circumcisi erant populus 5:5 Now all the people that came out were circumcised:

but all the people that were born in the wilderness by the way as they came forth out of Egypt, them they had not

autem qui natus est in deserto 5:6 per quadraginta annos circumcised.

5:6 For the children of Israel walked forty years in the wilderness, till all the people that were men of war, which came out of Egypt, were consumed, because they obeyed not the voice of the LORD: unto whom the LORD sware that he would not shew them the land, which the LORD sware unto their fathers that he would give us, a land that floweth with milk and honey.

itineris latissimae solitudinis incircumcisus fuit donec consumerentur qui non audierant uocem Domini et quibus ante iurauerat ut ostenderet eis terram lacte et melle manantem 5:7 horum filii in locum successerunt patrum et circumcisi 5:7 And their children, whom he raised up in their stead,

them Joshua circumcised: for they were uncircumcised, because they had not circumcised them by the way.

sunt ab Iosue quia sicut nati fuerant in praeputio erant nec eos in uia aliquis circumciderat 5:8 postquam autem om- 5:8 And it came to pass, when they had done circumcising

all the people, that they abode in their places in the camp, till they were whole.

nes circumcisi sunt manserunt in eodem castrorum loco donec sanarentur 5:9 dixitque Dominus ad Iosue hodie abs- 5:9 And the LORD said unto Joshua, This day have I

rolled away the reproach of Egypt from off you. Wherefore the name of the place is called Gilgal unto this day.

tuli obprobrium Aegypti a uobis uocatumque est nomen loci illius Galgala usque in praesentem diem 5:10 manserunt- 5:10 And the children of Israel encamped in Gilgal, and

kept the passover on the fourteenth day of the month at even in the plains of Jericho.

que filii Israhel in Galgalis et fecerunt phase quartadecima die mensis ad uesperum in campestribus Hiericho 5:11 et 5:11 And they did eat of the old corn of the land on the

morrow after the passover, unleavened cakes, and parched corn in the selfsame day.

comederunt de frugibus terrae die altero azymos panes et pulentam eiusdem anni 5:12 defecitque manna postquam 5:12 And the manna ceased on the morrow after they had

eaten of the old corn of the land; neither had the children of Israel manna any more; but they did eat of the fruit of

comederunt de frugibus terrae nec usi sunt ultra illo cibo fi- the land of Canaan that year. lii Israhel sed comederunt de frugibus praesentis anni terrae Chanaan 5:13 cum autem esset Iosue in agro urbis Hiericho 5:13 And it came to pass, when Joshua was by Jericho,

that he lifted up his eyes and looked, and, behold, there stood a man over against him with his sword drawn in his hand: and Joshua went unto him, and said unto him, Art thou for us, or for our adversaries?

leuauit oculos et uidit uirum stantem contra se et euaginatum tenentem gladium perrexitque ad eum et ait noster es an aduersariorum 5:14 qui respondit nequaquam sed sum 5:14 And he said, Nay; but as captain of the host of the

LORD am I now come.

princeps exercitus Domini et nunc uenio 5:15 cecidit Iosue 5:15 And Joshua fell on his face to the earth, and did wor-

ship, and said unto him, What saith my Lord unto his servant?

pronus in terram et adorans ait quid dominus meus loquitur

Biblia Sacra Vulgata

5:16 And the captain of the LORD’s host said unto Joshua, ad seruum suum 5:16 solue inquit calciamentum de pedi-Loose thy shoe from off thy foot; for the place whereon thou standest is holy. And Joshua did so.

bus tuis locus enim in quo stas sanctus est fecitque Iosue ut sibi fuerat imperatum

6:1 Now Jericho was straitly shut up because of the chil- 6:1 Hiericho autem clausa erat atque munita timore filio-

dren of Israel: none went out, and none came in.

6:2 And the LORD said unto Joshua, See, I have given into rum Israhel et nullus egredi audebat aut ingredi 6:2 dixit-

thine hand Jericho, and the king thereof, and the mighty men of valour.

que Dominus ad Iosue ecce dedi in manus tuas Hiericho et

6:3 And ye shall compass the city, all ye men of war, and regem eius omnesque fortes uiros 6:3 circuite urbem cuncti

go round about the city once. Thus shalt thou do six days. 6:4 And seven priests shall bear before the ark seven trum- bellatores semel per diem sic facietis sex diebus 6:4 seppets of rams’ horns: and the seventh day ye shall compass the city seven times, and the priests shall blow with the

trumpets. timo autem die sacerdotes tollant septem bucinas quarum

usus est in iobeleo et praecedant arcam foederis septies-

6:5 And it shall come to pass, that when they make a long que circuibitis ciuitatem et sacerdotes clangent bucinis 6:5

blast with the ram’s horn, and when ye hear the sound of the trumpet, all the people shall shout with a great shout; and the wall of the city shall fall down flat, and the people shall ascend up every man straight before him.

cumque insonuerit uox tubae longior atque concisior et in auribus uestris increpuerit conclamabit omnis populus uociferatione maxima et muri funditus corruent ciuitatis ingre-

6:6 And Joshua the son of Nun called the priests, and said dienturque singuli per locum contra quem steterint 6:6 uo-

unto them, Take up the ark of the covenant, and let seven priests bear seven trumpets of rams’ horns before the ark

of the LORD. cauit ergo Iosue filius Nun sacerdotes et dixit ad eos tol-

lite arcam foederis et septem alii sacerdotes tollant septem

6:7 And he said unto the people, Pass on, and compass the iobeleorum bucinas et incedant ante arcam Domini 6:7 ad

city, and let him that is armed pass on before the ark of the LORD.

populum quoque ait uadite et circuite ciuitatem armati prae-

6:8 And it came to pass, when Joshua had spoken unto the cedentes arcam Domini 6:8 cumque Iosue uerba finisset et

people, that the seven priests bearing the seven trumpets of rams’ horns passed on before the LORD, and blew with the trumpets: and the ark of the covenant of the LORD followed them.

septem sacerdotes septem bucinis clangerent ante arcam foe-

6:9 And the armed men went before the priests that blew deris Domini 6:9 omnisque praecederet armatus exercitus

with the trumpets, and the rereward came after the ark, the priests going on, and blowing with the trumpets.

reliquum uulgus arcam sequebatur ac bucinis omnia con-

6:10 And Joshua had commanded the people, saying, Ye crepabant 6:10 praeceperat autem Iosue populo dicens non

shall not shout, nor make any noise with your voice, neither shall any word proceed out of your mouth, until the

day I bid you shout; then shall ye shout. clamabitis nec audietur uox uestra neque ullus sermo ex ore

uestro egredietur donec ueniat dies in quo dicam uobis cla-

6:11 So the ark of the LORD compassed the city, going mate et uociferamini 6:11 circuiuit ergo arca Domini ciui-

about it once: and they came into the camp, and lodged in the camp.

tatem semel per diem et reuersa in castra mansit ibi 6:12

6:12 And Joshua rose early in the morning, and the priests took up the ark of the LORD.

igitur Iosue de nocte consurgente tulerunt sacerdotes arcam

6:13 And seven priests bearing seven trumpets of rams’ Domini 6:13 et septem ex eis septem bucinas quarum in io-

horns before the ark of the LORD went on continually, and blew with the trumpets: and the armed men went before them; but the rereward came after the ark of the LORD, the priests going on, and blowing with the trumpets.

beleis usus est praecedebantque arcam Domini ambulantes atque clangentes et armatus populus ibat ante eos uulgus au-

6:14 And the second day they compassed the city once, tem reliquum sequebatur arcam et bucinis personabat 6:14

and returned into the camp: so they did six days.

circumieruntque ciuitatem secundo die semel et reuersi sunt in castra sic fecerunt sex diebus 6:15 die autem septimo 6:15 And it came to pass on the seventh day, that they rose

early about the dawning of the day, and compassed the city after the same manner seven times: only on that day they

diluculo consurgentes circumierunt urbem sicut dispositum compassed the city seven times. erat septies 6:16 cumque septimo circuitu clangerent buci- 6:16 And it came to pass at the seventh time, when the

priests blew with the trumpets, Joshua said unto the people, Shout; for the LORD hath given you the city.

nis sacerdotes dixit Iosue ad omnem Israhel uociferamini tradidit enim uobis Dominus ciuitatem 6:17 sitque ciuitas 6:17 And the city shall be accursed, even it, and all that

are therein, to the LORD: only Rahab the harlot shall live, she and all that are with her in the house, because she hid

haec anathema et omnia quae in ea sunt Domino sola Raab the messengers that we sent. meretrix uiuat cum uniuersis qui cum ea in domo sunt abscondit enim nuntios quos direximus 6:18 uos autem cauete 6:18 And ye, in any wise keep yourselves from the ac-

cursed thing, lest ye make yourselves accursed, when ye take of the accursed thing, and make the camp of Israel a

ne de his quae praecepta sunt quippiam contingatis et sitis curse, and trouble it. praeuaricationis rei et omnia castra Israhel sub peccato sint atque turbentur 6:19 quicquid autem auri et argenti fuerit 6:19 But all the silver, and gold, and vessels of brass and

iron, are consecrated unto the LORD: they shall come into the treasury of the LORD.

et uasorum aeneorum ac ferri Domino consecretur repositum in thesauris eius 6:20 igitur omni uociferante populo et 6:20 So the people shouted when the priests blew with the

trumpets: and it came to pass, when the people heard the sound of the trumpet, and the people shouted with a great shout, that the wall fell down flat, so that the people went up into the city, every man straight before him, and they took the city.

clangentibus tubis postquam in aures multitudinis uox sonitusque increpuit muri ilico corruerunt et ascendit unusquisque per locum qui contra se erat ceperuntque ciuitatem 6:21 6:21 And they utterly destroyed all that was in the city,

both man and woman, young and old, and ox, and sheep, and ass, with the edge of the sword.

et interfecerunt omnia quae erant in ea a uiro usque ad mulierem ab infante usque ad senem boues quoque et oues et asinos in ore gladii percusserunt 6:22 duobus autem uiris qui 6:22 But Joshua had said unto the two men that had spied

out the country, Go into the harlot’s house, and bring out thence the woman, and all that she hath, as ye sware unto

exploratores missi fuerant dixit Iosue ingredimini domum her. mulieris meretricis et producite eam omniaque quae illius sunt sicut illi iuramento firmastis 6:23 ingressique iuuenes 6:23 And the young men that were spies went in, and

brought out Rahab, and her father, and her mother, and her brethren, and all that she had; and they brought out all

eduxerunt Raab et parentes eius fratres quoque et cunctam her kindred, and left them without the camp of Israel. supellectilem ac cognationem illius et extra castra Israhel manere fecerunt 6:24 urbem autem et omnia quae in ea sunt 6:24 And they burnt the city with fire, and all that was

therein: only the silver, and the gold, and the vessels of brass and of iron, they put into the treasury of the house of

succenderunt absque argento et auro et uasis aeneis ac ferro the LORD. quae in aerarium Domini consecrarunt 6:25 Raab uero me- 6:25 And Joshua saved Rahab the harlot alive, and her fa-

ther’s household, and all that she had; and she dwelleth in Israel even unto this day; because she hid the messengers,

retricem et domum patris eius atque omnia quae habebat fe- which Joshua sent to spy out Jericho. cit Iosue uiuere et habitauerunt in medio Israhel usque in praesentem diem eo quod absconderit nuntios quos miserat ut explorarent Hiericho in tempore illo inprecatus est Iosue dicens 6:26 maledictus uir coram Domino qui suscitauerit 6:26 And Joshua adjured them at that time, saying, Cursed

be the man before the LORD, that riseth up and buildeth this city Jericho: he shall lay the foundation thereof in his firstborn, and in his youngest son shall he set up the gates of it.

Biblia Sacra Vulgata

et aedificauerit ciuitatem Hiericho in primogenito suo fundamenta illius iaciat et in nouissimo liberorum ponat portas

6:27 So the LORD was with Joshua; and his fame was eius 6:27 fuit ergo Dominus cum Iosue et nomen eius in

noised throughout all the country.

omni terra uulgatum est

7:1 But the children of Israel committed a trespass in the 7:1 filii autem Israhel praeuaricati sunt mandatum et usur-

accursed thing: for Achan, the son of Carmi, the son of Zabdi, the son of Zerah, of the tribe of Judah, took of the accursed thing: and the anger of the LORD was kindled against the children of Israel.

pauerunt de anathemate nam Achan filius Charmi filii Zabdi filii Zare de tribu Iuda tulit aliquid de anathemate iratusque

7:2 And Joshua sent men from Jericho to Ai, which is be- est Dominus contra filios Israhel 7:2 cumque mitteret Io-

side Bethaven, on the east of Bethel, and spake unto them, saying, Go up and view the country. And the men went up

and viewed Ai. sue de Hiericho uiros contra Ahi quae est iuxta Bethauen ad

orientalem plagam oppidi Bethel dixit eis ascendite et explo-

7:3 And they returned to Joshua, and said unto him, Let rate terram qui praecepta conplentes explorauerunt Ahi 7:3

not all the people go up; but let about two or three thousand men go up and smite Ai; and make not all the people

to labour thither; for they are but few. et reuersi dixerunt ei non ascendat omnis populus sed duo

uel tria milia uirorum pergant et deleant ciuitatem quare om-

7:4 So there went up thither of the people about three thou- nis populus frustra uexatur contra hostes paucissimos 7:4

sand men: and they fled before the men of Ai.

ascenderunt ergo tria milia pugnatores qui statim terga uer-

7:5 And the men of Ai smote of them about thirty and six tentes 7:5 percussi sunt a uiris urbis Ahi et corruerunt ex

men: for they chased them from before the gate even unto Shebarim, and smote them in the going down: wherefore

the hearts of the people melted, and became as water. eis triginta et sex homines persecutique sunt eos aduersarii de porta usque Sabarim et ceciderunt per prona fugientes pertimuitque cor populi et instar aquae liquefactum est

7:6 And Joshua rent his clothes, and fell to the earth upon 7:6 Iosue uero scidit uestimenta sua et cecidit pronus in ter-

his face before the ark of the LORD until the eventide, he and the elders of Israel, and put dust upon their heads.

ram coram arca Domini usque ad uesperum tam ipse quam omnes senes Israhel miseruntque puluerem super capita sua

7:7 And Joshua said, Alas, O LORD God, wherefore hast 7:7 et dixit Iosue heu Domine Deus quid uoluisti transducere

thou at all brought this people over Jordan, to deliver us into the hand of the Amorites, to destroy us? would to God we had been content, and dwelt on the other side Jordan! populum istum Iordanem fluuium ut traderes nos in manus

Amorrei et perderes utinam ut coepimus mansissemus trans

7:8 O LORD, what shall I say, when Israel turneth their Iordanem 7:8 mi Domine Deus quid dicam uidens Israhebacks before their enemies!

7:9 For the Canaanites and all the inhabitants of the land lem hostibus suis terga uertentem 7:9 audient Chananei et

shall hear of it, and shall environ us round, and cut off our name from the earth: and what wilt thou do unto thy great

name? omnes habitatores terrae ac pariter conglobati circumdabunt

nos atque delebunt nomen nostrum de terra et quid facies

7:10 And the LORD said unto Joshua, Get thee up; where- magno nomini tuo 7:10 dixitque Dominus ad Iosue surge fore liest thou thus upon thy face?

7:11 Israel hath sinned, and they have also transgressed cur iaces pronus in terra 7:11 peccauit Israhel et praeuari-

my covenant which I commanded them: for they have even taken of the accursed thing, and have also stolen, and dissembled also, and they have put it even among their own stuff.

catus est pactum meum tuleruntque de anathemate et furati

7:12 Therefore the children of Israel could not stand be- sunt atque mentiti et absconderunt inter uasa sua 7:12 nec

fore their enemies, but turned their backs before their enemies, because they were accursed: neither will I be with you any more, except ye destroy the accursed from among you.

poterit Israhel stare ante hostes suos eosque fugiet quia pollutus est anathemate non ero ultra uobiscum donec conte-

ratis eum qui huius sceleris reus est 7:13 surge sanctifica 7:13 Up, sanctify the people, and say, Sanctify yourselves against to morrow: for thus saith the LORD God of Israel, There is an accursed thing in the midst of thee, O Israel: thou canst not stand before thine enemies, until ye take away the accursed thing from among you.

populum et dic eis sanctificamini in crastinum haec enim dicit Dominus Deus Israhel anathema in medio tui est Israhel non poteris stare coram hostibus tuis donec deleatur ex te qui hoc contaminatus est scelere 7:14 accedetisque 7:14 In the morning therefore ye shall be brought accord-

ing to your tribes: and it shall be, that the tribe which the LORD taketh shall come according to the families thereof; and the family which the LORD shall take shall come by households; and the household which the LORD shall take shall come man by man.

mane singuli per tribus uestras et quamcumque tribum sors inuenerit accedet per cognationes suas et cognatio per do-

mos domusque per uiros 7:15 et quicumque ille in hoc faci- 7:15 And it shall be, that he that is taken with the accursed thing shall be burnt with fire, he and all that he hath: because he hath transgressed the covenant of the LORD, and

nore fuerit deprehensus conburetur igni cum omni substantia because he hath wrought folly in Israel. sua quoniam praeuaricatus est pactum Domini et fecit nefas in Israhel 7:16 surgens itaque Iosue mane adplicauit Isra- 7:16 So Joshua rose up early in the morning, and brought

Israel by their tribes; and the tribe of Judah was taken:

hel per tribus suas et inuenta est tribus Iuda 7:17 quae cum 7:17 And he brought the family of Judah; and he took the

family of the Zarhites: and he brought the family of the Zarhites man by man; and Zabdi was taken:

iuxta familias suas esset oblata inuenta est familia Zarai illam quoque per uiros offerens repperit Zabdi 7:18 cuius do- 7:18 And he brought his household man by man; and

Achan, the son of Carmi, the son of Zabdi, the son of Zerah, of the tribe of Judah, was taken.

mum in singulos diuidens uiros inuenit Achan filium Charmi filii Zabdi filii Zare de tribu Iuda 7:19 et ait ad Achan fili 7:19 And Joshua said unto Achan, My son, give, I pray

thee, glory to the LORD God of Israel, and make confession unto him; and tell me now what thou hast done; hide

mi da gloriam Domino Deo Israhel et confitere atque indica it not from me. mihi quid feceris ne abscondas 7:20 responditque Achan 7:20 And Achan answered Joshua, and said, Indeed I have

sinned against the LORD God of Israel, and thus and thus have I done:

Iosue et dixit ei uere ego peccaui Domino Deo Israhel et sic et sic feci 7:21 uidi enim inter spolia pallium coccineum 7:21 When I saw among the spoils a goodly Babylonish

garment, and two hundred shekels of silver, and a wedge of gold of fifty shekels weight, then I coveted them, and took them; and, behold, they are hid in the earth in the midst of my tent, and the silver under it.

ualde bonum et ducentos siclos argenti regulamque auream quinquaginta siclorum et concupiscens abstuli et abscondi in terra contra medium tabernaculi mei argentumque fossa humo operui 7:22 misit ergo Iosue ministros qui currentes 7:22 So Joshua sent messengers, and they ran unto the

tent; and, behold, it was hid in his tent, and the silver under it.

ad tabernaculum illius reppererunt cuncta abscondita in eodem loco et argentum simul 7:23 auferentesque de tentorio 7:23 And they took them out of the midst of the tent, and

brought them unto Joshua, and unto all the children of Israel, and laid them out before the LORD.

tulerunt ea ad Iosue et ad omnes filios Israhel proieceruntque ante Dominum 7:24 tollens itaque Iosue Achan filium 7:24 And Joshua, and all Israel with him, took Achan the

son of Zerah, and the silver, and the garment, and the wedge of gold, and his sons, and his daughters, and his oxen, and his asses, and his sheep, and his tent, and all that he had: and they brought them unto the valley of Achor.

Zare argentumque et pallium et auream regulam filiosque eius et filias boues et asinos et oues ipsumque tabernaculum et cunctam supellectilem et omnis Israhel cum eo duxerunt

Biblia Sacra Vulgata

7:25 And Joshua said, Why hast thou troubled us? the eos ad uallem Achor 7:25 ubi dixit Iosue quia turbasti nos

LORD shall trouble thee this day. And all Israel stoned him with stones, and burned them with fire, after they had

stoned them with stones. exturbet te Dominus in die hac lapidauitque eum omnis Is-7:26 And they raised over him a great heap of stones unto rahel et cuncta quae illius erant igne consumpta sunt 7:26

this day. So the LORD turned from the fierceness of his anger. Wherefore the name of that place was called, The

valley of Achor, unto this day. congregauerunt quoque super eum aceruum magnum lapidum qui permanet usque in praesentem diem et auersus est furor Domini ab eis uocatumque est nomen loci illius uallis Achor usque hodie

8:1 And the LORD said unto Joshua, Fear not, neither be 8:1 dixit autem Dominus ad Iosue ne timeas neque formi-

thou dismayed: take all the people of war with thee, and arise, go up to Ai: see, I have given into thy hand the king

of Ai, and his people, and his city, and his land: des tolle tecum omnem multitudinem pugnatorum et consur-

gens ascende in oppidum Ahi ecce tradidi in manu tua regem

8:2 And thou shalt do to Ai and her king as thou didst unto eius et populum urbemque et terram 8:2 faciesque urbi Ahi

Jericho and her king: only the spoil thereof, and the cattle thereof, shall ye take for a prey unto yourselves: lay thee

an ambush for the city behind it. et regi eius sicut fecisti Hiericho et regi illius praedam uero

et omnia animantia diripietis uobis pone insidias urbi post

8:3 So Joshua arose, and all the people of war, to go up eam 8:3 surrexitque Iosue et omnis exercitus bellatorum

against Ai: and Joshua chose out thirty thousand mighty men of valour, and sent them away by night.

cum eo ut ascenderent in Ahi et electa triginta milia uiro-

8:4 And he commanded them, saying, Behold, ye shall lie rum fortium misit nocte 8:4 praecepitque eis dicens ponite

in wait against the city, even behind the city: go not very far from the city, but be ye all ready:

insidias post ciuitatem nec longius recedatis et eritis omnes

8:5 And I, and all the people that are with me, will ap- parati 8:5 ego autem et reliqua multitudo quae mecum est

proach unto the city: and it shall come to pass, when they come out against us, as at the first, that we will flee before

them, accedemus ex aduerso contra urbem cumque exierint con-8:6 (For they will come out after us) till we have drawn tra nos sicut ante fecimus fugiemus et terga uertemus 8:6

them from the city; for they will say, They flee before us, as at the first: therefore we will flee before them.

donec persequentes ab urbe longius protrahantur putabunt

8:7 Then ye shall rise up from the ambush, and seize upon enim fugere nos sicut prius 8:7 nobis ergo fugientibus et il-

the city: for the LORD your God will deliver it into your hand.

lis sequentibus consurgetis de insidiis et uastabitis ciuitatem

8:8 And it shall be, when ye have taken the city, that ye tradetque eam Dominus Deus uester in manus uestras 8:8

shall set the city on fire: according to the commandment of the LORD shall ye do. See, I have commanded you.

cumque ceperitis succendite eam sic omnia facietis ut iussi

8:9 Joshua therefore sent them forth: and they went to lie 8:9 dimisitque eos et perrexerunt ad insidiarum locum se-

in ambush, and abode between Bethel and Ai, on the west side of Ai: but Joshua lodged that night among the people.

deruntque inter Bethel et Ahi ad occidentalem plagam urbis

8:10 And Joshua rose up early in the morning, and num- Ahi Iosue autem nocte illa in medio mansit populi 8:10 sur-

bered the people, and went up, he and the elders of Israel, before the people to Ai.

gensque diluculo recensuit socios et ascendit cum senioribus

8:11 And all the people, even the people of war that were in fronte exercitus uallatus auxilio pugnatorum 8:11 cum-

with him, went up, and drew nigh, and came before the city, and pitched on the north side of Ai: now there was a

valley between them and Ai. que uenissent et ascendissent ex aduerso ciuitatis steterunt

ad septentrionalem urbis plagam inter quam et eos uallis me-

8:12 And he took about five thousand men, and set them dia erat 8:12 quinque milia autem uiros elegerat et posuerat

to lie in ambush between Bethel and Ai, on the west side of the city.

in insidiis inter Bethauen et Ahi ex occidentali parte eiusdem ciuitatis 8:13 omnis uero reliquus exercitus ad aquilonem 8:13 And when they had set the people, even all the host

that was on the north of the city, and their liers in wait on the west of the city, Joshua went that night into the midst

aciem dirigebat ita ut nouissimi multitudinis occidentalem of the valley. plagam urbis adtingerent abiit ergo Iosue nocte illa et stetit in uallis medio 8:14 quod cum uidisset rex Ahi festinauit 8:14 And it came to pass, when the king of Ai saw it, that

they hasted and rose up early, and the men of the city went out against Israel to battle, he and all his people, at a time appointed, before the plain; but he wist not that there were liers in ambush against him behind the city.

mane et egressus est cum omni exercitu ciuitatis direxitque aciem contra desertum ignorans quod post tergum laterent insidiae 8:15 Iosue uero et omnis Israhel cesserunt loco si- 8:15 And Joshua and all Israel made as if they were beaten

before them, and fled by the way of the wilderness.

mulantes metum et fugientes per uiam solitudinis 8:16 at 8:16 And all the people that were in Ai were called to-

gether to pursue after them: and they pursued after Joshua, and were drawn away from the city.

illi uociferantes pariter et se mutuo cohortantes persecuti sunt eos cumque recessissent a ciuitate 8:17 et ne unus qui- 8:17 And there was not a man left in Ai or Bethel, that

went not out after Israel: and they left the city open, and pursued after Israel.

dem in urbe Ahi et Bethel remansisset qui non persequeretur Israhel sicut eruperant aperta oppida relinquentes 8:18 dixit 8:18 And the LORD said unto Joshua, Stretch out the

spear that is in thy hand toward Ai; for I will give it into thine hand. And Joshua stretched out the spear that he had

Dominus ad Iosue leua clypeum qui in manu tua est contra in his hand toward the city.

urbem Ahi quoniam tibi tradam eam 8:19 cumque eleuas- 8:19 And the ambush arose quickly out of their place, and they ran as soon as he had stretched out his hand: and they entered into the city, and took it, and hasted and set the

set clypeum ex aduerso ciuitatis insidiae quae latebant sur- city on fire. rexerunt confestim et pergentes ad ciuitatem ceperunt et succenderunt eam 8:20 uiri autem ciuitatis qui persequebantur 8:20 And when the men of Ai looked behind them, they

saw, and, behold, the smoke of the city ascended up to heaven, and they had no power to flee this way or that way: and the people that fled to the wilderness turned back upon the pursuers.

Iosue respicientes et uidentes fumum urbis ad caelum usque conscendere non potuerunt ultra huc illucque diffugere praesertim cum hii qui simulauerant fugam et tendebant ad solitudinem contra persequentes fortissime restitissent 8:21 8:21 And when Joshua and all Israel saw that the ambush

had taken the city, and that the smoke of the city ascended, then they turned again, and slew the men of Ai.

uidensque Iosue et omnis Israhel quod capta esset ciuitas et fumus urbis ascenderet reuersus percussit uiros Ahi 8:22 si- 8:22 And the other issued out of the city against them; so

they were in the midst of Israel, some on this side, and some on that side: and they smote them, so that they let

quidem et illi qui ceperant et succenderant ciuitatem egressi none of them remain or escape. ex urbe contra suos medios hostium ferire coeperunt cum ergo ex utraque parte aduersarii caederentur ita ut nullus de tanta multitudine saluaretur 8:23 regem quoque urbis Ahi 8:23 And the king of Ai they took alive, and brought him

to Joshua.

adprehendere uiuentem et obtulerunt Iosue 8:24 igitur om- 8:24 And it came to pass, when Israel had made an end of

slaying all the inhabitants of Ai in the field, in the wilderness wherein they chased them, and when they were all fallen on the edge of the sword, until they were consumed, that all the Israelites returned unto Ai, and smote it with the edge of the sword.

nibus interfectis qui Israhelem ad deserta tendentem fuerant persecuti et in eodem loco gladio corruentibus reuersi filii Israhel percusserunt ciuitatem 8:25 erant autem qui in eo 8:25 And so it was, that all that fell that day, both of men

and women, were twelve thousand, even all the men of Ai.

die conciderant a uiro usque ad mulierem duodecim milia

Biblia Sacra Vulgata

8:26 For Joshua drew not his hand back, wherewith he hominum omnes urbis Ahi 8:26 Iosue uero non contraxit

stretched out the spear, until he had utterly destroyed all the inhabitants of Ai.

manum quam in sublime porrexerat tenens clypeum donec

8:27 Only the cattle and the spoil of that city Israel took interficerentur omnes habitatores Ahi 8:27 iumenta autem

for a prey unto themselves, according unto the word of the LORD which he commanded Joshua.

et praedam ciuitatis diuiserunt sibi filii Israhel sicut praece-

8:28 And Joshua burnt Ai, and made it an heap for ever, perat Dominus Iosue 8:28 qui succendit urbem et fecit eam

even a desolation unto this day. 8:29 And the king of Ai he hanged on a tree until eventide: tumulum sempiternum 8:29 regem quoque eius suspendit and as soon as the sun was down, Joshua commanded that they should take his carcase down from the tree, and cast it at the entering of the gate of the city, and raise thereon a great heap of stones, that remaineth unto this day.

in patibulo usque ad uesperum et solis occasum praecepitque et deposuerunt cadauer eius de cruce proieceruntque in ipso introitu ciuitatis congesto super eum magno aceruo lapidum

8:30 Then Joshua built an altar unto the LORD God of qui permanet usque in praesentem diem 8:30 tunc aedifi-

Israel in mount Ebal,

8:31 As Moses the servant of the LORD commanded the cauit Iosue altare Domino Deo Israhel in monte Hebal 8:31

children of Israel, as it is written in the book of the law of Moses, an altar of whole stones, over which no man hath lift up any iron: and they offered thereon burnt offerings unto the LORD, and sacrificed peace offerings.

sicut praeceperat Moses famulus Domini filiis Israhel et scriptum est in uolumine legis Mosi altare de lapidibus inpolitis quos ferrum non tetigit et obtulit super eo holocausta Do-

8:32 And he wrote there upon the stones a copy of the law mino immolauitque pacificas uictimas 8:32 et scripsit super

of Moses, which he wrote in the presence of the children of Israel.

lapides deuteronomium legis Mosi quod ille digesserat co-

8:33 And all Israel, and their elders, and officers, and their ram filiis Israhel 8:33 omnis autem populus et maiores natu judges, stood on this side the ark and on that side before the priests the Levites, which bare the ark of the covenant of the LORD, as well the stranger, as he that was born among them; half of them over against mount Gerizim, and half of them over against mount Ebal; as Moses the servant of the LORD had commanded before, that they should bless the people of Israel.

ducesque ac iudices stabant ex utraque parte arcae in conspectu sacerdotum qui portabant arcam foederis Domini ut aduena ita et indigena media eorum pars iuxta montem Garizim et media iuxta montem Hebal sicut praeceperat Moses famulus Domini et primum quidem benedixit populo Israhel

8:34 And afterward he read all the words of the law, the 8:34 post haec legit omnia uerba benedictionis et maledic-

blessings and cursings, according to all that is written in the book of the law.

tionis et cuncta quae scripta erant in legis uolumine 8:35

8:35 There was not a word of all that Moses commanded, which Joshua read not before all the congregation of Israel, with the women, and the little ones, and the strangers that were conversant among them.

nihil ex his quae Moses iusserat reliquit intactum sed uniuersa replicauit coram omni multitudine Israhel mulieribus ac paruulis et aduenis qui inter eos morabantur

9:1 And it came to pass, when all the kings which were 9:1 quibus auditis cuncti reges trans Iordanem qui uersa-

on this side Jordan, in the hills, and in the valleys, and in all the coasts of the great sea over against Lebanon, the Hittite, and the Amorite, the Canaanite, the Perizzite, the Hivite, and the Jebusite, heard thereof;

bantur in montanis et in campestribus in maritimis ac litore maris Magni hii quoque qui habitabant iuxta Libanum Hettheus et Amorreus et Chananeus Ferezeus et Eueus et Ie-

9:2 That they gathered themselves together, to fight with buseus 9:2 congregati sunt pariter ut pugnarent contra Io-

Joshua and with Israel, with one accord.

9:3 And when the inhabitants of Gibeon heard what sue et Israhel uno animo eademque sententia 9:3 at hii qui

Joshua had done unto Jericho and to Ai,

habitabant in Gabaon audientes cuncta quae fecerat Iosue Hiericho et Ahi 9:4 et callide cogitantes tulerunt sibi ciba- 9:4 They did work wilily, and went and made as if they had

been ambassadors, and took old sacks upon their asses, and wine bottles, old, and rent, and bound up;

ria saccos ueteres asinis inponentes et utres uinarios scissos atque consutos 9:5 calciamentaque perantiqua quae ad in- 9:5 And old shoes and clouted upon their feet, and old

garments upon them; and all the bread of their provision was dry and mouldy.

dicium uetustatis pittaciis consuta erant induti ueteribus uestimentis panes quoque quos portabant ob uiaticum duri erant et in frusta comminuti 9:6 perrexeruntque ad Iosue qui 9:6 And they went to Joshua unto the camp at Gilgal, and

said unto him, and to the men of Israel, We be come from a far country: now therefore make ye a league with us.

tunc morabatur in castris Galgalae et dixerunt ei atque omni simul Israheli de terra longinqua uenimus pacem uobiscum facere cupientes responderuntque uiri Israhel ad eos atque dixerunt 9:7 ne forsitan in terra quae nobis sorte debetur 9:7 And the men of Israel said unto the Hivites, Peradven-

ture ye dwell among us; and how shall we make a league with you?

habitetis et non possimus foedus inire uobiscum 9:8 at illi

9:8 And they said unto Joshua, We are thy servants. And Joshua said unto them, Who are ye? and from whence come ye?

ad Iosue serui inquiunt tui sumus quibus Iosue quinam ait estis et unde uenistis 9:9 responderunt de terra longinqua 9:9 And they said unto him, From a very far country thy

servants are come because of the name of the LORD thy God: for we have heard the fame of him, and all that he

ualde uenerunt serui tui in nomine Domini Dei tui audiui- did in Egypt, mus enim famam potentiae eius cuncta quae fecit in Aegypto 9:10 et duobus Amorreorum regibus trans Iordanem Seon re- 9:10 And all that he did to the two kings of the Amorites,

that were beyond Jordan, to Sihon king of Heshbon, and to Og king of Bashan, which was at Ashtaroth.

gi Esebon et Og regi Basan qui erat in Astharoth 9:11 dixe-

9:11 Wherefore our elders and all the inhabitants of our country spake to us, saying, Take victuals with you for the journey, and go to meet them, and say unto them, We are your servants: therefore now make ye a league with us.

runtque nobis seniores et omnes habitatores terrae nostrae tollite in manibus cibaria ob longissimam uiam et occurrite eis ac dicite serui uestri sumus foedus inite nobiscum 9:12 9:12 This our bread we took hot for our provision out of

our houses on the day we came forth to go unto you; but now, behold, it is dry, and it is mouldy:

en panes quando egressi sumus de domibus nostris ut ueniremus ad uos calidos sumpsimus nunc sicci facti sunt et uetustate nimia comminuti 9:13 utres uini nouos impleui- 9:13 And these bottles of wine, which we filled, were new;

and, behold, they be rent: and these our garments and our shoes are become old by reason of the very long journey.

mus nunc rupti sunt et soluti uestes et calciamenta quibus induimur et quae habemus in pedibus ob longitudinem largioris uiae trita sunt et paene consumpta 9:14 susceperunt 9:14 And the men took of their victuals, and asked not

counsel at the mouth of the LORD.

igitur de cibariis eorum et os Domini non interrogauerunt 9:15 fecitque Iosue cum eis pacem et inito foedere pollici- 9:15 And Joshua made peace with them, and made a

league with them, to let them live: and the princes of the congregation sware unto them.

tus est quod non occiderentur principes quoque multitudinis iurauerunt eis 9:16 post dies autem tres initi foederis audie- 9:16 And it came to pass at the end of three days after they

had made a league with them, that they heard that they were their neighbours, and that they dwelt among them.

runt quod in uicino habitarent et inter eos futuri essent 9:17

9:17 And the children of Israel journeyed, and came unto their cities on the third day. Now their cities were Gibeon, and Chephirah, and Beeroth, and Kirjathjearim.

moueruntque castra filii Israhel et uenerunt in ciuitates eo-

Biblia Sacra Vulgata

rum die tertio quarum haec uocabula sunt Gabaon et Caphira

9:18 And the children of Israel smote them not, because et Beroth et Cariathiarim 9:18 et non percusserunt eos eo

the princes of the congregation had sworn unto them by the LORD God of Israel. And all the congregation mur-

mured against the princes. quod iurassent eis principes multitudinis in nomine Domini

Dei Israhel murmurauit itaque omne uulgus contra principes

9:19 But all the princes said unto all the congregation, We 9:19 qui responderunt eis iurauimus illis in nomine Domini

have sworn unto them by the LORD God of Israel: now therefore we may not touch them.

Dei Israhel et idcirco non possumus eos contingere 9:20

9:20 This we will do to them; we will even let them live, lest wrath be upon us, because of the oath which we sware unto them.

sed hoc faciemus eis reseruentur quidem ut uiuant ne con-

9:21 And the princes said unto them, Let them live; but let tra nos ira Domini concitetur si peierauerimus 9:21 sed sic

them be hewers of wood and drawers of water unto all the congregation; as the princes had promised them.

uiuant ut in usus uniuersae multitudinis ligna caedant aquas-

9:22 And Joshua called for them, and he spake unto them, que conportent quibus haec loquentibus 9:22 uocauit Ga-

saying, Wherefore have ye beguiled us, saying, We are very far from you; when ye dwell among us?

baonitas Iosue et dixit eis cur nos decipere fraude uoluistis ut diceretis procul ualde habitamus a uobis cum in medio

9:23 Now therefore ye are cursed, and there shall none of nostri sitis 9:23 itaque sub maledictione eritis et non defi-

you be freed from being bondmen, and hewers of wood and drawers of water for the house of my God.

ciet de stirpe uestra ligna caedens aquasque conportans in

9:24 And they answered Joshua, and said, Because it was domum Dei mei 9:24 qui responderunt nuntiatum est nobis

certainly told thy servants, how that the LORD thy God commanded his servant Moses to give you all the land, and to destroy all the inhabitants of the land from before you, therefore we were sore afraid of our lives because of you, and have done this thing.

seruis tuis quae promisisset Dominus Deus tuus Mosi seruo suo ut traderet uobis omnem terram et disperderet cunctos habitatores eius timuimus igitur ualde et prouidimus animabus nostris uestro terrore conpulsi et hoc consilium iniuimus

9:25 And now, behold, we are in thine hand: as it seemeth 9:25 nunc autem in manu tua sumus quod tibi bonum et regood and right unto thee to do unto us, do.

9:26 And so did he unto them, and delivered them out of ctum uidetur fac nobis 9:26 fecit ergo Iosue ut dixerat et

the hand of the children of Israel, that they slew them not.

liberauit eos de manibus filiorum Israhel ut non occiderentur

9:27 And Joshua made them that day hewers of wood and 9:27 decreuitque in illo die esse eos in ministerium cuncti po-

drawers of water for the congregation, and for the altar of the LORD, even unto this day, in the place which he

should choose. puli et altaris Domini caedentes ligna et aquas conportantes usque in praesens tempus in loco quem Dominus elegisset

10:1 Now it came to pass, when Adonizedec king of 10:1 quae cum audisset Adonisedec rex Hierusalem quod

Jerusalem had heard how Joshua had taken Ai, and had utterly destroyed it; as he had done to Jericho and her king, so he had done to Ai and her king; and how the inhabitants of Gibeon had made peace with Israel, and were among them;

scilicet cepisset Iosue Ahi et subuertisset eam sicut enim fecerat Hiericho et regi eius sic fecit Ahi et regi illius et quod transfugissent Gabaonitae ad Israhel et essent foederati eo-

10:2 That they feared greatly, because Gibeon was a great rum 10:2 timuit ualde urbs enim magna erat Gabaon et una city, as one of the royal cities, and because it was greater than Ai, and all the men thereof were mighty.

regalium ciuitatum et maior oppido Ahi omnesque bellato-

10:3 Wherefore Adonizedec king of Jerusalem, sent unto res eius fortissimi 10:3 misit ergo Adonisedec rex Hierusa-

Hoham king of Hebron, and unto Piram king of Jarmuth, and unto Japhia king of Lachish, and unto Debir king of

Eglon, saying, lem ad Oham regem Hebron et ad Pharam regem Hieremoth ad Iaphie quoque regem Lachis et ad Dabir regem Eglon dicens 10:4 ascendite ad me et ferte praesidium ut expug- 10:4 Come up unto me, and help me, that we may smite

Gibeon: for it hath made peace with Joshua and with the children of Israel.

nemus Gabaon quare transfugerit ad Iosue et filios Israhel 10:5 congregati igitur ascenderunt quinque reges Amorreo- 10:5 Therefore the five kings of the Amorites, the king of

Jerusalem, the king of Hebron, the king of Jarmuth, the king of Lachish, the king of Eglon, gathered themselves together, and went up, they and all their hosts, and encamped before Gibeon, and made war against it.

rum rex Hierusalem rex Hebron rex Hieremoth rex Lachis rex Eglon simul cum exercitibus suis et castrametati sunt circa Gabaon obpugnantes eam 10:6 habitatores autem Ga- 10:6 And the men of Gibeon sent unto Joshua to the camp

to Gilgal, saying, Slack not thy hand from thy servants; come up to us quickly, and save us, and help us: for all the kings of the Amorites that dwell in the mountains are gathered together against us.

baon urbis obsessae miserunt ad Iosue qui tunc morabatur in castris apud Galgalam et dixerunt ei ne retrahas manus tuas ab auxilio seruorum tuorum ascende cito et libera nos ferque praesidium conuenerunt enim aduersum nos omnes reges Amorreorum qui habitant in montanis 10:7 ascenditque 10:7 So Joshua ascended from Gilgal, he, and all the peo-

ple of war with him, and all the mighty men of valour.

Iosue de Galgalis et omnis exercitus bellatorum cum eo uiri fortissimi 10:8 dixitque Dominus ad Iosue ne timeas eos in 10:8 And the LORD said unto Joshua, Fear them not: for I

have delivered them into thine hand; there shall not a man of them stand before thee.

manus enim tuas tradidi illos nullus tibi ex eis resistere poterit 10:9 inruit itaque Iosue super eos repente tota ascendens 10:9 Joshua therefore came unto them suddenly, and went

up from Gilgal all night.

nocte de Galgalis 10:10 et conturbauit eos Dominus a fa- 10:10 And the LORD discomfited them before Israel, and

slew them with a great slaughter at Gibeon, and chased them along the way that goeth up to Bethhoron, and smote

cie Israhel contriuitque plaga magna in Gabaon ac persecu- them to Azekah, and unto Makkedah. tus est per uiam ascensus Bethoron et percussit usque Azeca et Maceda 10:11 cumque fugerent filios Israhel et essent in 10:11 And it came to pass, as they fled from before Is-

rael, and were in the going down to Bethhoron, that the LORD cast down great stones from heaven upon them unto Azekah, and they died: they were more which died with hailstones than they whom the children of Israel slew with the sword.

descensu Bethoron Dominus misit super eos lapides magnos de caelo usque Azeca et mortui sunt multo plures lapidibus grandinis quam quos gladio percusserant filii Israhel 10:12 10:12 Then spake Joshua to the LORD in the day when

the LORD delivered up the Amorites before the children of Israel, and he said in the sight of Israel, Sun, stand thou

tunc locutus est Iosue Domino in die qua tradidit Amorreum still upon Gibeon; and thou, Moon, in the valley of Ajalon. in conspectu filiorum Israhel dixitque coram eis sol contra

Gabaon ne mouearis et luna contra uallem Ahialon 10:13 10:13 And the sun stood still, and the moon stayed, until the people had avenged themselves upon their enemies. Is not this written in the book of Jasher? So the sun stood still in the midst of heaven, and hasted not to go down about a whole day.

steteruntque sol et luna donec ulcisceretur se gens de inimicis suis nonne scriptum est hoc in libro Iustorum stetit itaque sol in medio caeli et non festinauit occumbere spatio unius diei 10:14 non fuit ante et postea tam longa dies oboediente 10:14 And there was no day like that before it or after it,

that the LORD hearkened unto the voice of a man: for the LORD fought for Israel.

Domino uoci hominis et pugnante pro Israhel 10:15 reuer-

10:15 And Joshua returned, and all Israel with him, unto the camp to Gilgal.

susque est Iosue cum omni Israhel in castra Galgalae 10:16 _{10:16 But these five kings fled, and hid themselves in a}

cave at Makkedah.

fugerant enim quinque reges et se absconderant in spelunca

Biblia Sacra Vulgata

10:17 And it was told Joshua, saying, The five kings are urbis Maceda 10:17 nuntiatumque est Iosue quod inuenti

found hid in a cave at Makkedah.

10:18 And Joshua said, Roll great stones upon the mouth essent quinque reges latentes in spelunca Maceda 10:18 qui

of the cave, and set men by it for to keep them:

praecepit sociis et ait uoluite saxa ingentia ad os spelun-

10:19 And stay ye not, but pursue after your enemies, and cae et ponite uiros industrios qui clausos custodiant 10:19

smite the hindmost of them; suffer them not to enter into their cities: for the LORD your God hath delivered them

into your hand. uos autem nolite stare sed persequimini hostes et extremos quosque fugientium caedite ne dimittatis eos urbium suarum intrare praesidia quos tradidit Dominus Deus in manus ue-

10:20 And it came to pass, when Joshua and the children stras 10:20 caesis igitur aduersariis plaga magna et usque

of Israel had made an end of slaying them with a very great slaughter, till they were consumed, that the rest which re-

mained of them entered into fenced cities. ad internicionem paene consumptis hii qui Israhel effugere 10:21 And all the people returned to the camp to Joshua potuerunt ingressi sunt ciuitates munitas 10:21 reuersusque

at Makkedah in peace: none moved his tongue against any of the children of Israel.

est omnis exercitus ad Iosue in Maceda ubi tunc erant castra sani et integro numero nullusque contra filios Israhel muttire

10:22 Then said Joshua, Open the mouth of the cave, and ausus est 10:22 praecepitque Iosue dicens aperite os spe-

bring out those five kings unto me out of the cave.

luncae et producite ad me quinque reges qui in ea latitant

10:23 And they did so, and brought forth those five kings 10:23 fecerunt ministri ut sibi fuerat imperatum et eduxerunt

unto him out of the cave, the king of Jerusalem, the king of Hebron, the king of Jarmuth, the king of Lachish, and

the king of Eglon. ad eum quinque reges de spelunca regem Hierusalem regem 10:24 And it came to pass, when they brought out those Hebron regem Hieremoth regem Lachis regem Eglon 10:24

kings unto Joshua, that Joshua called for all the men of Israel, and said unto the captains of the men of war which went with him, Come near, put your feet upon the necks of these kings. And they came near, and put their feet upon the necks of them.

cumque educti essent ad eum uocauit omnes uiros Israhel et ait ad principes exercitus qui secum erant ite et ponite pedes super colla regum istorum qui cum perrexissent et subiecto-

10:25 And Joshua said unto them, Fear not, nor be dis- rum pedibus colla calcarent 10:25 rursum ait ad eos nolite

mayed, be strong and of good courage: for thus shall the LORD do to all your enemies against whom ye fight.

timere nec paueatis confortamini et estote robusti sic enim faciet Dominus cunctis hostibus uestris aduersum quos di-

10:26 And afterward Joshua smote them, and slew them, micatis 10:26 percussitque Iosue et interfecit eos atque su-

and hanged them on five trees: and they were hanging upon the trees until the evening.

spendit super quinque stipites fueruntque suspensi usque ad

10:27 And it came to pass at the time of the going down of uesperum 10:27 cumque occumberet sol praecepit sociis ut

the sun, that Joshua commanded, and they took them down off the trees, and cast them into the cave wherein they had been hid, and laid great stones in the cave’s mouth, which remain until this very day.

deponerent eos de patibulis qui depositos proiecerunt in speluncam in qua latuerant et posuerunt super os eius saxa in-

10:28 And that day Joshua took Makkedah, and smote it gentia quae permanent usque in praesens 10:28 eodem die

with the edge of the sword, and the king thereof he utterly destroyed, them, and all the souls that were therein; he let none remain: and he did to the king of Makkedah as he did unto the king of Jericho.

Macedam quoque cepit Iosue et percussit in ore gladii regemque illius interfecit et omnes habitatores eius non dimisit in ea saltim paruas reliquias fecitque regi Maceda sicut

10:29 Then Joshua passed from Makkedah, and all Israel fecerat regi Hiericho 10:29 transiuit cum omni Israhel de

with him, unto Libnah, and fought against Libnah:

Maceda in Lebna et pugnabat contra eam 10:30 quam tra- 10:30 And the LORD delivered it also, and the king

thereof, into the hand of Israel; and he smote it with the edge of the sword, and all the souls that were therein; he let none remain in it; but did unto the king thereof as he did unto the king of Jericho.

didit Dominus cum rege suo in manu Israhel percusseruntque urbem in ore gladii et omnes habitatores eius non dimiserunt in ea ullas reliquias feceruntque regi Lebna sicut fecerant regi Hiericho 10:31 de Lebna transiuit in Lachis 10:31 And Joshua passed from Libnah, and all Israel with

him, unto Lachish, and encamped against it, and fought against it:

et exercitu per gyrum disposito obpugnabat eam 10:32 tra-

10:32 And the LORD delivered Lachish into the hand of Israel, which took it on the second day, and smote it with the edge of the sword, and all the souls that were therein, according to all that he had done to Libnah.

diditque Dominus Lachis in manu Israhel et cepit eam die altero atque percussit in ore gladii omnemque animam quae fuerat in ea sicut fecerat Lebna 10:33 eo tempore ascendit 10:33 Then Horam king of Gezer came up to help Lachish;

and Joshua smote him and his people, until he had left him none remaining.

Hiram rex Gazer ut auxiliaretur Lachis quem percussit Io-

sue cum omni populo eius usque ad internicionem 10:34 10:34 And from Lachish Joshua passed unto Eglon, and all Israel with him; and they encamped against it, and fought against it:

transiuitque de Lachis in Eglon et circumdedit 10:35 atque

10:35 And they took it on that day, and smote it with the edge of the sword, and all the souls that were therein he utterly destroyed that day, according to all that he had done to Lachish.

expugnauit eam eadem die percussitque in ore gladii omnes animas quae erant in ea iuxta omnia quae fecerat Lachis 10:36 ascendit quoque cum omni Israhele de Eglon in Heb- 10:36 And Joshua went up from Eglon, and all Israel with

him, unto Hebron; and they fought against it:

ron et pugnauit contra eam 10:37 cepitque et percussit in 10:37 And they took it, and smote it with the edge of the

sword, and the king thereof, and all the cities thereof, and all the souls that were therein; he left none remaining, according to all that he had done to Eglon; but destroyed it utterly, and all the souls that were therein.

ore gladii regem quoque eius et omnia oppida regionis illius uniuersasque animas quae in ea fuerant commoratae non reliquit in ea ullas reliquias sicut fecerat Eglon sic fecit et Hebron cuncta quae in ea repperit consumens gladio 10:38 inde 10:38 And Joshua returned, and all Israel with him, to De-

bir;

reuersus in Dabir 10:39 cepit eam atque uastauit 10:40 re- 10:39 and fought against it: And he took it,

10:40 and the king thereof, and all the cities thereof; and they smote them with the edge of the sword, and utterly destroyed all the souls that were therein; he left none remaining: as he had done to Hebron, so he did to Debir, and to the king thereof; as he had done also to Libnah, and to her king.

gem quoque eius et omnia per circuitum oppida percussit in ore gladii non dimisit in ea ullas reliquias sicut fecerat Hebron et Lebna et regibus earum sic fecit Dabir et regi illius 10:41 percussit itaque Iosue omnem terram montanam 10:41 So Joshua smote all the country of the hills, and of

the south, and of the vale, and of the springs, and all their kings: he left none remaining, but utterly destroyed all that

et meridianam atque campestrem et Asedoth cum regibus breathed, as the LORD God of Israel commanded. suis non dimisit in ea ullas reliquias sed omne quod spirare poterat interfecit sicut praeceperat ei Dominus Deus Israhel 10:42 a Cadesbarne usque Gazam omnem terram Gosen us- 10:42 And Joshua smote them from Kadeshbarnea even

unto Gaza, and all the country of Goshen, even unto Gibeon.

que Gabaon 10:43 uniuersos reges et regiones eorum uno

10:43 And all these kings and their land did Joshua take at one time, because the LORD God of Israel fought for Israel.

cepit impetu atque uastauit Dominus enim Deus Israhel pugnabat pro eo 10:44 reuersusque est cum omni Israhele ad 10:44 And Joshua returned, and all Israel with him, unto

the camp to Gilgal.

locum castrorum in Galgala

Biblia Sacra Vulgata

11:1 And it came to pass, when Jabin king of Hazor had 11:1 quae cum audisset Iabin rex Asor misit ad Iobab re-

heard those things, that he sent to Jobab king of Madon, and to the king of Shimron, and to the king of Achshaph,

gem Madon et ad regem Someron atque ad regem Acsaph

11:2 And to the kings that were on the north of the moun- 11:2 ad reges quoque aquilonis qui habitabant in montanis et

tains, and of the plains south of Chinneroth, and in the valley, and in the borders of Dor on the west,

in planitie contra meridiem Cheneroth in campestribus quo-

11:3 And to the Canaanite on the east and on the west, and que et in regionibus Dor iuxta mare 11:3 Chananeumque ab

to the Amorite, and the Hittite, and the Perizzite, and the Jebusite in the mountains, and to the Hivite under Hermon

in the land of Mizpeh. oriente et occidente et Amorreum atque Hettheum ac Fere-

zeum et Iebuseum in montanis Eueum quoque qui habitabat

11:4 And they went out, they and all their hosts with them, ad radices Hermon in terra Masphe 11:4 egressique sunt much people, even as the sand that is upon the sea shore in multitude, with horses and chariots very many.

omnes cum turmis suis populus multus nimis sicut harena quae est in litore maris equi quoque et currus inmensae mul-

11:5 And when all these kings were met together, they titudinis 11:5 conueneruntque omnes reges isti in unum ad

came and pitched together at the waters of Merom, to fight against Israel.

aquas Merom ut pugnarent contra Israhel 11:6 dixitque Do-

11:6 And the LORD said unto Joshua, Be not afraid because of them: for to morrow about this time will I deliver them up all slain before Israel: thou shalt hough their horses, and burn their chariots with fire.

minus ad Iosue ne timeas eos cras enim hac eadem hora ego tradam omnes istos uulnerandos in conspectu Israhel equos

11:7 So Joshua came, and all the people of war with him, eorum subneruabis et currus igne conbures 11:7 uenitque

against them by the waters of Merom suddenly; and they fell upon them.

Iosue et omnis exercitus cum eo aduersum illos ad aquas

11:8 And the LORD delivered them into the hand of Is- Merom subito et inruerunt super eos 11:8 tradiditque illos

rael, who smote them, and chased them unto great Zidon, and unto Misrephothmaim, and unto the valley of Mizpeh eastward; and they smote them, until they left them none remaining.

Dominus in manu Israhel qui percusserunt eos et persecuti sunt usque ad Sidonem magnam et aquas Maserefoth campumque Masphe qui est ad orientalem illius partem ita per-

11:9 And Joshua did unto them as the LORD bade him: cussit omnes ut nullas dimitteret ex eis reliquias 11:9 fe-

he houghed their horses, and burnt their chariots with fire.

cit sicut praeceperat ei Dominus equos eorum subneruauit

11:10 And Joshua at that time turned back, and took Ha- currusque conbusit 11:10 reuersusque statim cepit Asor et

zor, and smote the king thereof with the sword: for Hazor beforetime was the head of all those kingdoms.

regem eius percussit gladio Asor enim antiquitus inter om-

11:11 And they smote all the souls that were therein with nia regna haec principatum tenebat 11:11 percussitque om-

the edge of the sword, utterly destroying them: there was not any left to breathe: and he burnt Hazor with fire.

nes animas quae ibidem morabantur non dimisit in ea ullas reliquias sed usque ad internicionem uniuersa uastauit

11:12 And all the cities of those kings, and all the kings ipsamque urbem permisit incendio 11:12 et omnes per cir-

of them, did Joshua take, and smote them with the edge of the sword, and he utterly destroyed them, as Moses the

servant of the LORD commanded. cuitum ciuitates regesque earum cepit percussit atque dele-

11:13 But as for the cities that stood still in their strength, uit sicut praeceperat ei Moses famulus Domini 11:13 abs-Israel burned none of them, save Hazor only; that did Joshua burn.

que urbibus quae erant in collibus et in tumulis sitae ceteras succendit Israhel unam tantum Asor munitissimam flamma

11:14 And all the spoil of these cities, and the cattle, the consumpsit 11:14 omnemque praedam istarum urbium ac children of Israel took for a prey unto themselves; but every man they smote with the edge of the sword, until they had destroyed them, neither left they any to breathe.

iumenta diuiserunt sibi filii Israhel cunctis hominibus interfectis 11:15 sicut praeceperat Dominus Mosi seruo suo ita 11:15 As the LORD commanded Moses his servant, so

did Moses command Joshua, and so did Joshua; he left nothing undone of all that the LORD commanded Moses.

praecepit Moses Iosue et ille uniuersa conpleuit non praeteriit de uniuersis mandatis ne unum quidem uerbum quod iusserat Dominus Mosi 11:16 cepit itaque Iosue omnem ter- 11:16 So Joshua took all that land, the hills, and all the

south country, and all the land of Goshen, and the valley, and the plain, and the mountain of Israel, and the valley of

ram montanam et meridianam terramque Gosen et planitiem the same; et occidentalem plagam montemque Israhel et campestria eius 11:17 et partem montis quae ascendit Seir usque Baal- 11:17 Even from the mount Halak, that goeth up to Seir,

even unto Baalgad in the valley of Lebanon under mount Hermon: and all their kings he took, and smote them, and

gad per planitiem Libani subter montem Hermon omnes re- slew them. ges eorum cepit percussit occidit 11:18 multo tempore pu- 11:18 Joshua made war a long time with all those kings. gnauit Iosue contra reges istos 11:19 non fuit ciuitas quae 11:19 There was not a city that made peace with the chil-

dren of Israel, save the Hivites the inhabitants of Gibeon: all other they took in battle.

se non traderet filiis Israhel praeter Eueum qui habitabat in Gabaon omnes bellando cepit 11:20 Domini enim senten- 11:20 For it was of the LORD to harden their hearts, that

they should come against Israel in battle, that he might destroy them utterly, and that they might have no favour, but that he might destroy them, as the LORD commanded Moses.

tiae fuerat ut indurarentur corda eorum et pugnarent contra Israhel et caderent et non mererentur ullam clementiam ac perirent sicut praeceperat Dominus Mosi 11:21 in tempore 11:21 And at that time came Joshua, and cut off the

Anakims from the mountains, from Hebron, from Debir, from Anab, and from all the mountains of Judah, and from all the mountains of Israel: Joshua destroyed them utterly with their cities.

illo uenit Iosue et interfecit Enacim de montanis Hebron et Dabir et Anab et de omni monte Iuda et Israhel urbesque eorum deleuit 11:22 non reliquit ullum de stirpe Enacim 11:22 There was none of the Anakims left in the land of

the children of Israel: only in Gaza, in Gath, and in Ashdod, there remained.

in terra filiorum Israhel absque ciuitatibus Gaza et Geth et Azoto in quibus solis relicti sunt 11:23 cepit ergo Iosue om- 11:23 So Joshua took the whole land, according to all that

the LORD said unto Moses; and Joshua gave it for an inheritance unto Israel according to their divisions by their

nem terram sicut locutus est Dominus ad Mosen et tradidit tribes. And the land rested from war. eam in possessionem filiis Israhel secundum partes et tribus suas quieuitque terra a proeliis

12:1 hii sunt reges quos percusserunt filii Israhel et pos- 12:1 Now these are the kings of the land, which the chil-

dren of Israel smote, and possessed their land on the other side Jordan toward the rising of the sun, from the river

sederunt terram eorum trans Iordanem ad solis ortum a tor- Arnon unto mount Hermon, and all the plain on the east: rente Arnon usque ad montem Hermon et omnem orientalem plagam quae respicit solitudinem 12:2 Seon rex Amorreo- 12:2 Sihon king of the Amorites, who dwelt in Heshbon,

and ruled from Aroer, which is upon the bank of the river Arnon, and from the middle of the river, and from half Gilead, even unto the river Jabbok, which is the border of the children of Ammon;

rum qui habitauit in Esebon dominatus est ab Aroer quae sita est super ripam torrentis Arnon et mediae partis in ualle dimidiique Galaad usque ad torrentem Iaboc qui est terminus filiorum Ammon 12:3 et a solitudine usque ad mare Chene- 12:3 And from the plain to the sea of Chinneroth on the

east, and unto the sea of the plain, even the salt sea on the east, the way to Bethjeshimoth; and from the south, under

roth contra orientem et usque ad mare Deserti quod est mare Ashdothpisgah:

Salsissimum ad orientalem plagam per uiam quae ducit Bethesimoth et ab australi parte quae subiacent Asedothphasga

12:4 And the coast of Og king of Bashan, which was of 12:4 terminus Og regis Basan de reliquiis Rafaim qui habita-

the remnant of the giants, that dwelt at Ashtaroth and at Edrei,

uit in Astharoth et in Edrain et dominatus est in monte Hermon et in Salacha atque in uniuersa Basan usque ad terminos

12:5 And reigned in mount Hermon, and in Salcah, and 12:5 Gesuri et Machathi et dimidiae partis Galaad terminos

in all Bashan, unto the border of the Geshurites and the Maachathites, and half Gilead, the border of Sihon king of

Heshbon. Seon regis Esebon 12:6 Moses famulus Domini et filii Is-

12:6 Them did Moses the servant of the LORD and the children of Israel smite: and Moses the servant of the LORD gave it for a possession unto the Reubenites, and the Gadites, and the half tribe of Manasseh.

rahel percusserunt eos tradiditque terram eorum Moses in possessionem Rubenitis et Gadditis et dimidiae tribui Ma-

12:7 And these are the kings of the country which Joshua nasse 12:7 hii sunt reges terrae quos percussit Iosue et filii

and the children of Israel smote on this side Jordan on the west, from Baalgad in the valley of Lebanon even unto the mount Halak, that goeth up to Seir; which Joshua gave unto the tribes of Israel for a possession according to their divisions;

Israhel trans Iordanem ad occidentalem plagam a Baalgad in campo Libani usque ad montem cuius pars ascendit in Seir tradiditque eam Iosue in possessionem tribubus Israhel

12:8 In the mountains, and in the valleys, and in the plains, singulis partes suas 12:8 tam in montanis quam in planis and in the springs, and in the wilderness, and in the south country; the Hittites, the Amorites, and the Canaanites, the

Perizzites, the Hivites, and the Jebusites: atque campestribus in Aseroth et solitudine ac meridie Hettheus fuit et Amorreus Chananeus et Ferezeus Eueus et Ie-

12:9 The king of Jericho, one; the king of Ai, which is buseus 12:9 rex Hiericho unus rex Ahi quae est ex latere

beside Bethel, one;

12:10 The king of Jerusalem, one; the king of Hebron, Bethel unus 12:10 rex Hierusalem unus rex Hebron unus

one;

12:11 The king of Jarmuth, one; the king of Lachish, one; 12:11 rex Hierimoth unus rex Lachis unus 12:12 rex Eglon

12:12 The king of Eglon, one; the king of Gezer, one;

12:13 The king of Debir, one; the king of Geder, one; unus rex Gazer unus 12:13 rex Dabir unus rex Gader unus 12:14 The king of Hormah, one; the king of Arad, one; 12:14 rex Herma unus rex Hered unus 12:15 rex Lebna unus

12:15 The king of Libnah, one; the king of Adullam, one;

12:16 The king of Makkedah, one; the king of Bethel, one; rex Odollam unus 12:16 rex Maceda unus rex Bethel unus

12:17 The king of Tappuah, one; the king of Hepher, one; 12:17 rex Thaffua unus rex Afer unus 12:18 rex Afec unus

12:18 The king of Aphek, one; the king of Lasharon, one;

12:19 The king of Madon, one; the king of Hazor, one; rex Saron unus 12:19 rex Madon unus rex Asor unus 12:20

12:20 The king of Shimronmeron, one; the king of Achshaph, one;

rex Someron unus rex Acsaph unus 12:21 rex Thenach unus

12:21 The king of Taanach, one; the king of Megiddo, one;

12:22 The king of Kedesh, one; the king of Jokneam of rex Mageddo unus 12:22 rex Cades unus rex Iachanaem Carmel, one;

12:23 The king of Dor in the coast of Dor, one; the king Chermeli unus 12:23 rex Dor et prouinciae Dor unus rex

of the nations of Gilgal, one;

12:24 The king of Tirzah, one: all the kings thirty and one. gentium Galgal unus 12:24 rex Thersa unus omnes reges

triginta et unus

13:1 Now Joshua was old and stricken in years; and the 13:1 Iosue senex prouectaeque aetatis erat et dixit Domi-

LORD said unto him, Thou art old and stricken in years, and there remaineth yet very much land to be possessed.

nus ad eum senuisti et longeuus es terraque latissima dere-

13:2 This is the land that yet remaineth: all the borders of licta est quae necdum est sorte diuisa 13:2 omnis uidelithe Philistines, and all Geshuri,

13:3 From Sihor, which is before Egypt, even unto the bor- cet Galilea Philisthim et uniuersa Gesuri 13:3 a fluuio tur-

ders of Ekron northward, which is counted to the Canaanite: five lords of the Philistines; the Gazathites, and the Ashdothites, the Eshkalonites, the Gittites, and the Ekronites; also the Avites:

bido qui inrigat Aegyptum usque ad terminos Accaron contra aquilonem terra Chanaan quae in quinque regulos Philisthim diuiditur Gazeos Azotios Ascalonitas Gettheos et Accaronitas 13:4 ad meridiem uero sunt Euei omnis terra Cha- 13:4 From the south, all the land of the Canaanites, and

Mearah that is beside the Sidonians unto Aphek, to the borders of the Amorites:

naan et Maara Sidoniorum usque Afeca et terminos Amorrei 13:5 eiusque confinia Libani quoque regio contra orien- 13:5 And the land of the Giblites, and all Lebanon, toward

the sunrising, from Baalgad under mount Hermon unto the entering into Hamath.

tem a Baalgad sub monte Hermon donec ingrediaris Emath

13:6 omnium qui habitant in monte a Libano usque ad aquas 13:6 All the inhabitants of the hill country from Lebanon unto Misrephothmaim, and all the Sidonians, them will I drive out from before the children of Israel: only divide thou it by lot unto the Israelites for an inheritance, as I have commanded thee.

Masrefoth uniuersique Sidonii ego sum qui delebo eos a facie filiorum Israhel ueniat ergo in parte hereditatis Israhel sicut praecepi tibi 13:7 et nunc diuide terram in possessio- 13:7 Now therefore divide this land for an inheritance unto

the nine tribes, and the half tribe of Manasseh,

nem nouem tribubus et dimidiae tribui Manasse 13:8 cum 13:8 With whom the Reubenites and the Gadites have re-

ceived their inheritance, which Moses gave them, beyond Jordan eastward, even as Moses the servant of the LORD

qua Ruben et Gad possederunt terram quam tradidit eis Mo- gave them; ses famulus Domini trans fluenta Iordanis ad orientalem plagam 13:9 ab Aroer quae sita est in ripa torrentis Arnon et in 13:9 From Aroer, that is upon the bank of the river Arnon,

and the city that is in the midst of the river, and all the plain of Medeba unto Dibon;

uallis medio uniuersaque campestria Medaba usque Dibon 13:10 et cunctas ciuitates Seon regis Amorrei qui regnauit in 13:10 And all the cities of Sihon king of the Amorites,

which reigned in Heshbon, unto the border of the children of Ammon;

Esebon usque ad terminos filiorum Ammon 13:11 et Ga-

13:11 And Gilead, and the border of the Geshurites and Maachathites, and all mount Hermon, and all Bashan unto Salcah;

laad ac terminum Gesuri et Machathi omnemque montem Hermon et uniuersam Basan usque Saleca 13:12 omne re- 13:12 All the kingdom of Og in Bashan, which reigned in

Ashtaroth and in Edrei, who remained of the remnant of the giants: for these did Moses smite, and cast them out.

gnum Og in Basan qui regnauit in Astharoth et Edraim ipse fuit de reliquiis Rafaim percussitque eos Moses atque deleuit 13:13 nolueruntque disperdere filii Israhel Gesuri et Machathi 13:13 Nevertheless the children of Israel expelled not the

Geshurites, nor the Maachathites: but the Geshurites and the Maachathites dwell among the Israelites until this day.

et habitauerunt in medio Israhel usque in praesentem diem 13:14 tribui autem Leui non dedit possessionem sed sacrifi- 13:14 Only unto the tribes of Levi he gave none inheri-

tance; the sacrifices of the LORD God of Israel made by fire are their inheritance, as he said unto them.

cia et uictimae Domini Dei Israhel ipsa est eius hereditas sicut locutus est illi 13:15 dedit ergo Moses possessionem 13:15 And Moses gave unto the tribe of the children of

Reuben inheritance according to their families.

tribui filiorum Ruben iuxta cognationes suas 13:16 fuitque 13:16 And their coast was from Aroer, that is on the bank

of the river Arnon, and the city that is in the midst of the river, and all the plain by Medeba;

terminus eorum ab Aroer quae sita est in ripa torrentis Arnon et in ualle eiusdem torrentis media uniuersam planitiem

quae ducit Medaba 13:17 et Esebon cunctosque uiculos ea- 13:17 Heshbon, and all her cities that are in the plain; Di-

bon, and Bamothbaal, and Bethbaalmeon,

rum qui sunt in campestribus Dibon quoque et Bamothbaal et oppidum Baalmaon 13:18 Iessa et Cedmoth et Mephe- 13:18 And Jahaza, and Kedemoth, and Mephaath,

Biblia Sacra Vulgata

13:19 And Kirjathaim, and Sibmah, and Zarethshahar in eth 13:19 Cariathaim et Sebama et Sarathasar in monte con-

the mount of the valley,

13:20 And Bethpeor, and Ashdothpisgah, and Bethjeshi- uallis 13:20 Bethpheor et Asedothphasga et Bethaisimoth

moth,

13:21 And all the cities of the plain, and all the kingdom 13:21 omnes urbes campestres uniuersaque regna Seon regis

of Sihon king of the Amorites, which reigned in Heshbon, whom Moses smote with the princes of Midian, Evi, and Rekem, and Zur, and Hur, and Reba, which were dukes of Sihon, dwelling in the country.

Amorrei qui regnauit in Esebon quem percussit Moses cum principibus Madian Eueum et Recem et Sur et Vr et Rabee

13:22 Balaam also the son of Beor, the soothsayer, did duces Seon habitatores terrae 13:22 et Balaam filium Beor

the children of Israel slay with the sword among them that were slain by them.

ariolum occiderunt filii Israhel gladio cum ceteris interfectis

13:23 And the border of the children of Reuben was Jor- 13:23 factusque est terminus filiorum Ruben Iordanis fluuius

dan, and the border thereof. This was the inheritance of the children of Reuben after their families, the cities and

the villages thereof. haec est possessio Rubenitarum per cognationes suas urbium 13:24 And Moses gave inheritance unto the tribe of Gad, et uiculorum 13:24 deditque Moses tribui Gad et filiis eius

even unto the children of Gad according to their families.

per cognationes suas possessionem cuius haec diuisio est

13:25 And their coast was Jazer, and all the cities of 13:25 terminus Iazer et omnes ciuitates Galaad dimidiamque

Gilead, and half the land of the children of Ammon, unto Aroer that is before Rabbah;

partem terrae filiorum Ammon usque ad Aroer quae est con-

13:26 And from Heshbon unto Ramathmizpeh, and Be- tra Rabba 13:26 et ab Esebon usque Ramoth Masphe et Ba-

tonim; and from Mahanaim unto the border of Debir; 13:27 And in the valley, Betharam, and Bethnimrah, and tanim et a Manaim usque ad terminos Dabir 13:27 in ualle Succoth, and Zaphon, the rest of the kingdom of Sihon king of Heshbon, Jordan and his border, even unto the edge of the sea of Chinnereth on the other side Jordan eastward.

quoque Betharaam et Bethnemra et Soccoth et Saphon reliquam partem regni Seon regis Esebon huius quoque Iordanis finis est usque ad extremam partem maris Chenereth trans

13:28 This is the inheritance of the children of Gad after Iordanem ad orientalem plagam 13:28 haec est possessio fi-

their families, the cities, and their villages.

13:29 And Moses gave inheritance unto the half tribe of liorum Gad per familias suas ciuitates et uillae earum 13:29

Manasseh: and this was the possession of the half tribe of the children of Manasseh by their families.

dedit et dimidiae tribui Manasse filiisque eius iuxta cogna-

13:30 And their coast was from Mahanaim, all Bashan, all tiones suas possessionem 13:30 cuius hoc principium est

the kingdom of Og king of Bashan, and all the towns of Jair, which are in Bashan, threescore cities:

a Manaim uniuersam Basan et cuncta regna Og regis Basan omnesque uicos Air qui sunt in Basan sexaginta oppida

13:31 And half Gilead, and Ashtaroth, and Edrei, cities of 13:31 et dimidiam partem Galaad Astharoth et Edrai urbes the kingdom of Og in Bashan, were pertaining unto the children of Machir the son of Manasseh, even to the one

half of the children of Machir by their families. regni Og in Basan filiis Machir filii Manasse dimidiae parti 13:32 These are the countries which Moses did distribute filiorum Machir iuxta cognationes suas 13:32 hanc posses-

for inheritance in the plains of Moab, on the other side Jordan, by Jericho, eastward.

sionem diuisit Moses in campestribus Moab trans Iordanem

13:33 But unto the tribe of Levi Moses gave not any in- contra Hiericho ad orientalem plagam 13:33 tribui autem

heritance: the LORD God of Israel was their inheritance, as he said unto them.

Leui non dedit possessionem quoniam Dominus Deus Israhel ipse est possessio eius ut locutus est illi

14:1 And these are the countries which the children of Is- 14:1 hoc est quod possederunt filii Israhel in terra Charael inherited in the land of Canaan, which Eleazar the priest, and Joshua the son of Nun, and the heads of the fathers of the tribes of the children of Israel, distributed for inheritance to them.

naan quam dederunt eis Eleazar sacerdos et Iosue filius Nun et principes familiarum per tribus Israhel 14:2 sorte omnia 14:2 By lot was their inheritance, as the LORD com-

manded by the hand of Moses, for the nine tribes, and for the half tribe.

diuidentes sicut praeceperat Dominus in manu Mosi nouem tribubus et dimidiae tribui 14:3 duabus enim tribubus et di- 14:3 For Moses had given the inheritance of two tribes and

an half tribe on the other side Jordan: but unto the Levites he gave none inheritance among them.

midiae dederat Moses trans Iordanem possessionem absque Leuitis qui nihil terrae acceperunt inter fratres suos 14:4 14:4 For the children of Joseph were two tribes, Manasseh

and Ephraim: therefore they gave no part unto the Levites in the land, save cities to dwell in, with their suburbs for

sed in eorum successerant locum filii Ioseph in duas diuisi their cattle and for their substance. tribus Manasse et Ephraim nec acceperunt Leuitae aliam in terra partem nisi urbes ad habitandum et suburbana earum ad alenda iumenta et pecora sua 14:5 sicut praecepit 14:5 As the LORD commanded Moses, so the children of

Israel did, and they divided the land.

Dominus Mosi ita fecerunt filii Israhel et diuiserunt terram 14:6 accesserunt itaque filii Iuda ad Iosue in Galgala locu- 14:6 Then the children of Judah came unto Joshua in Gil-

gal: and Caleb the son of Jephunneh the Kenezite said unto him, Thou knowest the thing that the LORD said unto Moses the man of God concerning me and thee in Kadeshbarnea.

tusque est ad eum Chaleb filius Iepphonne Cenezeus nosti quid locutus sit Dominus ad Mosen hominem Dei de me et te in Cadesbarne 14:7 quadraginta annorum eram quando 14:7 Forty years old was I when Moses the servant of the

LORD sent me from Kadeshbarnea to espy out the land; and I brought him word again as it was in mine heart.

me misit Moses famulus Domini de Cadesbarne ut considerarem terram nuntiauique ei quod mihi uerum uidebatur 14:8 fratres autem mei qui ascenderant mecum dissoluerunt 14:8 Nevertheless my brethren that went up with me made

the heart of the people melt: but I wholly followed the LORD my God.

cor populi et nihilominus ego secutus sum Dominum De-

um meum 14:9 iurauitque Moses in die illo dicens terram 14:9 And Moses sware on that day, saying, Surely the land whereon thy feet have trodden shall be thine inheritance, and thy children’s for ever, because thou hast wholly fol-

quam calcauit pes tuus erit possessio tua et filiorum tuorum lowed the LORD my God.

in aeternum quia secutus es Dominum Deum meum 14:10 14:10 And now, behold, the LORD hath kept me alive, as he said, these forty and five years, even since the LORD spake this word unto Moses, while the children of Israel wandered in the wilderness: and now, lo, I am this day fourscore and five years old.

concessit ergo Dominus uitam mihi sicut pollicitus est usque in praesentem diem quadraginta et quinque anni sunt ex quo locutus est Dominus uerbum istud ad Mosen quando ambulabat Israhel per solitudinem hodie octoginta quinque annorum sum 14:11 sic ualens ut eo ualebam tempore quando ad 14:11 As yet I am as strong this day as I was in the day

that Moses sent me: as my strength was then, even so is my strength now, for war, both to go out, and to come in.

explorandum missus sum illius in me temporis fortitudo usque hodie perseuerat tam ad bellandum quam ad gradiendum 14:12 da ergo mihi montem istum quem pollicitus est Domi- 14:12 Now therefore give me this mountain, whereof the

LORD spake in that day; for thou heardest in that day how the Anakims were there, and that the cities were great and fenced: if so be the LORD will be with me, then I shall be able to drive them out, as the LORD said.

nus te quoque audiente in quo Enacim sunt et urbes magnae atque munitae si forte sit Dominus mecum et potuero delere eos sicut promisit mihi 14:13 benedixitque ei Iosue et tradi- 14:13 And Joshua blessed him, and gave unto Caleb the

son of Jephunneh Hebron for an inheritance.

dit Hebron in possessionem 14:14 atque ex eo fuit Hebron 14:14 Hebron therefore became the inheritance of Caleb

the son of Jephunneh the Kenezite unto this day, because that he wholly followed the LORD God of Israel.

Biblia Sacra Vulgata

Chaleb filio Iepphonne Cenezeo usque in praesentem diem

14:15 And the name of Hebron before was Kirjatharba; quia secutus est Dominum Deum Israhel 14:15 nomen He-

which Arba was a great man among the Anakims. And the land had rest from war.

bron antea uocabatur Cariatharbe Adam maximus ibi inter Enacim situs est et terra cessauit a proeliis

15:1 This then was the lot of the tribe of the children of 15:1 igitur sors filiorum Iudae per cognationes suas ista

Judah by their families; even to the border of Edom the wilderness of Zin southward was the uttermost part of the

south coast. fuit a termino Edom desertum Sin contra meridiem et us-15:2 And their south border was from the shore of the salt que ad extremam partem australis plagae 15:2 initium eius

sea, from the bay that looketh southward:

a summitate maris Salsissimi et a lingua eius quae respicit

15:3 And it went out to the south side to Maalehacrabbim, meridiem 15:3 egrediturque contra ascensum Scorpionis et

and passed along to Zin, and ascended up on the south side unto Kadeshbarnea, and passed along to Hezron, and went

up to Adar, and fetched a compass to Karkaa: pertransit in Sina ascenditque in Cadesbarne et peruenit in 15:4 From thence it passed toward Azmon, and went out Esrom ascendens Addara et circumiens Caricaa 15:4 atque

unto the river of Egypt; and the goings out of that coast were at the sea: this shall be your south coast.

inde pertransiens in Asemona et perueniens ad torrentem Aegypti eruntque termini eius mare Magnum hic erit finis

15:5 And the east border was the salt sea, even unto the meridianae plagae 15:5 ab oriente uero erit initium mare

end of Jordan. And their border in the north quarter was from the bay of the sea at the uttermost part of Jordan:

Salsissimum usque ad extrema Iordanis et ea quae respiciunt aquilonem a lingua maris usque ad eundem Iordanem

15:6 And the border went up to Bethhogla, and passed fluuium 15:6 ascenditque terminus in Bethagla et transit

along by the north of Betharabah; and the border went up to the stone of Bohan the son of Reuben:

ab aquilone in Betharaba ascendens ad lapidem Boem filii

15:7 And the border went up toward Debir from the valley Ruben 15:7 et tendens usque ad terminos Debera de ualle of Achor, and so northward, looking toward Gilgal, that is before the going up to Adummim, which is on the south side of the river: and the border passed toward the waters of Enshemesh, and the goings out thereof were at Enrogel:

Achor contra aquilonem respiciens Galgala quae est ex aduerso ascensionis Adommim ab australi parte torrentis transitque aquas quae uocantur fons Solis et erunt exitus eius

15:8 And the border went up by the valley of the son ad fontem Rogel 15:8 ascenditque per conuallem filii En-

of Hinnom unto the south side of the Jebusite; the same is Jerusalem: and the border went up to the top of the mountain that lieth before the valley of Hinnom westward, which is at the end of the valley of the giants northward:

nom ex latere Iebusei ad meridiem haec est Hierusalem et inde se erigens ad uerticem montis qui est contra Gehennom ad occidentem in summitate uallis Rafaim contra aquilonem

15:9 And the border was drawn from the top of the hill 15:9 pertransitque a uertice montis usque ad fontem aquae

unto the fountain of the water of Nephtoah, and went out to the cities of mount Ephron; and the border was drawn

to Baalah, which is Kirjathjearim: Nepthoa et peruenit usque ad uicos montis Ephron inclina-

turque in Bala quae est Cariathiarim id est urbs Siluarum

15:10 And the border compassed from Baalah westward 15:10 et circuit de Bala contra occidentem usque ad mon-

unto mount Seir, and passed along unto the side of mount Jearim, which is Chesalon, on the north side, and went

down to Bethshemesh, and passed on to Timnah: tem Seir transitque iuxta latus montis Iarim ad aquilonem

in Cheslon et descendit in Bethsames transitque in Thamna

15:11 And the border went out unto the side of Ekron 15:11 et peruenit contra aquilonem partis Accaron ex latere

northward: and the border was drawn to Shicron, and passed along to mount Baalah, and went out unto Jabneel; and the goings out of the border were at the sea.

inclinaturque Sechrona et transit montem Baala peruenitque in Iebnehel et maris Magni contra occidentem fine concluditur 15:12 hii sunt termini filiorum Iuda per circuitum in 15:12 And the west border was to the great sea, and the

coast thereof. This is the coast of the children of Judah round about according to their families.

cognationibus suis 15:13 Chaleb uero filio Iepphonne dedit

15:13 And unto Caleb the son of Jephunneh he gave a part among the children of Judah, according to the commandment of the LORD to Joshua, even the city of Arba the father of Anak, which city is Hebron.

partem in medio filiorum Iuda sicut praeceperat ei Dominus Cariatharbe patris Enach ipsa est Hebron 15:14 deleuitque 15:14 And Caleb drove thence the three sons of Anak,

Sheshai, and Ahiman, and Talmai, the children of Anak.

ex ea Chaleb tres filios Enach Sesai et Ahiman et Tholmai de stirpe Enach 15:15 atque inde conscendens uenit ad habita- 15:15 And he went up thence to the inhabitants of Debir:

and the name of Debir before was Kirjathsepher.

tores Dabir quae prius uocabatur Cariathsepher id est ciuitas Litterarum 15:16 dixitque Chaleb qui percusserit Cariath- 15:16 And Caleb said, He that smiteth Kirjathsepher, and

taketh it, to him will I give Achsah my daughter to wife.

sepher et ceperit eam dabo illi Axam filiam meam uxorem 15:17 cepitque eam Othonihel filius Cenez frater Chaleb iu- 15:17 And Othniel the son of Kenaz, the brother of Caleb,

took it: and he gave him Achsah his daughter to wife.

nior deditque ei Axam filiam suam uxorem 15:18 quae cum 15:18 And it came to pass, as she came unto him, that she

moved him to ask of her father a field: and she lighted off her ass; and Caleb said unto her, What wouldest thou?

pergerent simul suasit uiro ut peteret a patre suo agrum suspirauitque ut sedebat in asino cui Chaleb quid habes inquit 15:19 at illa respondit da mihi benedictionem terram austra- 15:19 Who answered, Give me a blessing; for thou hast

given me a south land; give me also springs of water. And he gave her the upper springs, and the nether springs.

lem et arentem dedisti mihi iunge et inriguam dedit itaque ei Chaleb inriguum superius et inferius 15:20 haec est posses- 15:20 This is the inheritance of the tribe of the children of

Judah according to their families.

sio tribus filiorum Iuda per cognationes suas 15:21 erantque 15:21 And the uttermost cities of the tribe of the chil-

dren of Judah toward the coast of Edom southward were Kabzeel, and Eder, and Jagur,

ciuitates ab extremis partibus filiorum Iuda iuxta terminos Edom a meridie Cabsehel et Eder et Iagur 15:22 et Cina et 15:22 And Kinah, and Dimonah, and Adadah, Dimona Adeda 15:23 et Cedes et Asor Iethnan 15:24 Zif et 15:23 And Kedesh, and Hazor, and Ithnan,

15:24 Ziph, and Telem, and Bealoth,

Thelem Baloth 15:25 et Asor noua et Cariothesrom haec est _{15:25 And Hazor, Hadattah, and Kerioth, and Hezron,}

which is Hazor,

Asor 15:26 Aman Same et Molada 15:27 et Asergadda et 15:26 Amam, and Shema, and Moladah,

15:27 And Hazargaddah, and Heshmon, and Bethpalet,

Asemon Bethfeleth 15:28 et Asersual et Bersabee et Bazio- _{15:28 And Hazarshual, and Beersheba, and Bizjothjah,} thia 15:29 Bala et Hiim Esem 15:30 et Heltholad Exiil et 15:29 Baalah, and Iim, and Azem,

15:30 And Eltolad, and Chesil, and Hormah,

Harma 15:31 Siceleg et Medemena et Sensenna 15:32 Le- _{15:31 And Ziklag, and Madmannah, and Sansannah,}

15:32 And Lebaoth, and Shilhim, and Ain, and Rimmon: all the cities are twenty and nine, with their villages:

baoth et Selim et Aenremmon omnes ciuitates uiginti nouem et uillae earum 15:33 in campestribus uero Esthaul et Saraa 15:33 And in the valley, Eshtaol, and Zoreah, and Ashnah, et Asena 15:34 et Azanoe et Aengannim Thaffua et Aenaim 15:34 And Zanoah, and Engannim, Tappuah, and Enam, 15:35 et Hierimoth Adulam Soccho et Azeca 15:36 et Sa- 15:35 Jarmuth, and Adullam, Socoh, and Azekah,

15:36 And Sharaim, and Adithaim, and Gederah, and Gederothaim; fourteen cities with their villages:

raim Adithaim et Gedera et Giderothaim urbes quattuordecim et uillae earum 15:37 Sanan et Adesa et Magdalgad 15:37 Zenan, and Hadashah, and Migdalgad,

Biblia Sacra Vulgata

15:38 And Dilean, and Mizpeh, and Joktheel, 15:38 Delean et Mesfa et Iecthel 15:39 Lachis et Bascath et

15:39 Lachish, and Bozkath, and Eglon,

15:40 And Cabbon, and Lahmam, and Kithlish, Aglon 15:40 Thebbon et Lehemas et Chethlis 15:41 et Gi-

15:41 And Gederoth, Bethdagon, and Naamah, and Makkedah; sixteen cities with their villages:

deroth Bethdagon et Neema et Maceda ciuitates sedecim et

15:42 Libnah, and Ether, and Ashan, uillae earum 15:42 Labana et Aether et Asan 15:43 Ieptha

15:43 And Jiphtah, and Ashnah, and Nezib,

15:44 And Keilah, and Achzib, and Mareshah; nine cities et Esna et Nesib 15:44 Ceila et Achzib et Maresa ciuitates

with their villages:

15:45 Ekron, with her towns and her villages: nouem et uillae earum 15:45 Accaron cum uicis et uillulis

15:46 From Ekron even unto the sea, all that lay near Ash- suis 15:46 ab Accaron usque ad mare omnia quae uergunt dod, with their villages:

15:47 Ashdod with her towns and her villages, Gaza with ad Azotum et uiculos eius 15:47 Azotus cum uicis et uillu-

her towns and her villages, unto the river of Egypt, and the great sea, and the border thereof:

lis suis Gaza cum uiculis et uillulis suis usque ad torrentem

15:48 And in the mountains, Shamir, and Jattir, and Socoh, Aegypti mare Magnum terminus eius 15:48 et in monte Sa-

15:49 And Dannah, and Kirjathsannah, which is Debir, mir et Iether et Soccho 15:49 et Edenna Cariathsenna haec 15:50 And Anab, and Eshtemoh, and Anim, est Dabir 15:50 Anab et Isthemo et Anim 15:51 Gosen et

15:51 And Goshen, and Holon, and Giloh; eleven cities with their villages:

Olon et Gilo ciuitates undecim et uillae earum 15:52 Arab

15:52 Arab, and Dumah, and Eshean,

15:53 And Janum, and Bethtappuah, and Aphekah, et Roma et Esaan 15:53 Ianum et Bethafua et Afeca 15:54

15:54 And Humtah, and Kirjatharba, which is Hebron, and Zior; nine cities with their villages:

Ammatha et Cariatharbe haec est Hebron et Sior ciuitates

15:55 Maon, Carmel, and Ziph, and Juttah, nouem et uillae earum 15:55 Maon et Chermel et Zif et Io-15:56 And Jezreel, and Jokdeam, and Zanoah, tae 15:56 Iezrehel et Iucadam et Zanoe 15:57 Accaim Ge-

15:57 Cain, Gibeah, and Timnah; ten cities with their villages:

baa et Thamna ciuitates decem et uillae earum 15:58 Alul

15:58 Halhul, Bethzur, and Gedor, 15:59 And Maarath, and Bethanoth, and Eltekon; six cities et Bethsur et Gedor 15:59 Mareth et Bethanoth et Elthewith their villages:

15:60 Kirjathbaal, which is Kirjathjearim, and Rabbah; cen ciuitates sex et uillae earum 15:60 Cariathbaal haec est

two cities with their villages:

Cariathiarim urbs Siluarum et Arebba ciuitates duae et uil-

15:61 In the wilderness, Betharabah, Middin, and Se- lae earum 15:61 in deserto Betharaba Meddin et Schacha

cacah,

15:62 And Nibshan, and the city of Salt, and Engedi; six 15:62 Anepsan et ciuitas Salis et Engaddi ciuitates sex et uil-

cities with their villages. 15:63 As for the Jebusites the inhabitants of Jerusalem, lae earum 15:63 Iebuseum autem habitatorem Hierusalem the children of Judah could not drive them out; but the Jebusites dwell with the children of Judah at Jerusalem

unto this day. non potuerunt filii Iuda delere habitauitque Iebuseus cum filiis Iuda in Hierusalem usque in praesentem diem

16:1 And the lot of the children of Joseph fell from Jordan 16:1 cecidit quoque sors filiorum Ioseph ab Iordane conby Jericho, unto the water of Jericho on the east, to the wilderness that goeth up from Jericho throughout mount

Bethel, tra Hiericho et aquas eius ab oriente solitudo quae ascendit 16:2 And goeth out from Bethel to Luz, and passeth along de Hiericho ad montana Bethel 16:2 et egreditur de Bethel

unto the borders of Archi to Ataroth,

16:3 And goeth down westward to the coast of Japhleti, Luzam transitque terminum Archiatharoth 16:3 et descen-

unto the coast of Bethhoron the nether, and to Gezer; and the goings out thereof are at the sea.

dit ad occidentem iuxta terminum Ieflethi usque ad terminos Bethoron inferioris et Gazer finiunturque regiones eius

16:4 So the children of Joseph, Manasseh and Ephraim, mari Magno 16:4 possederuntque filii Ioseph Manasse et

took their inheritance.

Ephraim 16:5 et factus est terminus filiorum Ephraim per 16:5 And the border of the children of Ephraim according

to their families was thus: even the border of their inheritance on the east side was Atarothaddar, unto Bethhoron

cognationes suas et possessio eorum contra orientem Atha- the upper; rothaddar usque Bethoron superiorem 16:6 egrediunturque 16:6 And the border went out toward the sea to Mich-

methah on the north side; and the border went about eastward unto Taanathshiloh, and passed by it on the east to

confinia in mare Machmethath uero aquilonem respicit et Janohah; circuit terminus contra orientem in Thanathselo et pertransit ab oriente Ianoe 16:7 descenditque de Ianoe in Atharoth 16:7 And it went down from Janohah to Ataroth, and to

Naarath, and came to Jericho, and went out at Jordan.

et Noaratha et peruenit in Hiericho et egreditur ad Iordanem 16:8 de Taffua pertransitque contra mare in ualle Harundineti 16:8 The border went out from Tappuah westward unto

the river Kanah; and the goings out thereof were at the sea. This is the inheritance of the tribe of the children of

suntque egressus eius in mare Salsissimum haec est posses- Ephraim by their families. sio tribus filiorum Ephraim per familias suas 16:9 urbesque 16:9 And the separate cities for the children of Ephraim

were among the inheritance of the children of Manasseh, all the cities with their villages.

quae separatae sunt filiis Ephraim in medio possessionis filiorum Manasse et uillae earum 16:10 et non interfecerunt 16:10 And they drave not out the Canaanites that dwelt in

Gezer: but the Canaanites dwell among the Ephraimites unto this day, and serve under tribute.

filii Ephraim Chananeum qui habitabat in Gazer habitauitque Chananeus in medio Ephraim usque in diem hanc tributarius

17:1 cecidit autem sors tribui Manasse ipse est enim pri- 17:1 There was also a lot for the tribe of Manasseh; for he

was the firstborn of Joseph; to wit, for Machir the firstborn of Manasseh, the father of Gilead: because he was a man

mogenitus Ioseph Machir primogenito Manasse patri Galaad of war, therefore he had Gilead and Bashan. qui fuit uir pugnator habuitque possessionem Galaad et Basan 17:2 et reliquis filiorum Manasse iuxta familias suas 17:2 There was also a lot for the rest of the children of

Manasseh by their families; for the children of Abiezer, and for the children of Helek, and for the children of Asriel, and for the children of Shechem, and for the children of Hepher, and for the children of Shemida: these were the male children of Manasseh the son of Joseph by their families.

filiis Abiezer et filiis Elech et filiis Esrihel et filiis Sechem et filiis Epher et filiis Semida isti sunt filii Manasse filii Ioseph mares per cognationes suas 17:3 Salphaad uero filio Epher 17:3 But Zelophehad, the son of Hepher, the son of Gilead,

the son of Machir, the son of Manasseh, had no sons, but daughters: and these are the names of his daughters,

filii Galaad filii Machir filii Manasse non erant filii sed solae Mahlah, and Noah, Hoglah, Milcah, and Tirzah. filiae quarum ista sunt nomina Maala et Noa Egla et Melcha et Thersa 17:4 ueneruntque in conspectu Eleazari sacerdo- 17:4 And they came near before Eleazar the priest, and

before Joshua the son of Nun, and before the princes, saying, The LORD commanded Moses to give us an inheritance among our brethren. Therefore according to the commandment of the LORD he gave them an inheritance among the brethren of their father.

tis et Iosue filii Nun et principum dicentes Dominus praecepit per manum Mosi ut daretur nobis possessio in medio fratrum nostrorum deditque eis iuxta imperium Domini possessionem in medio fratrum patris earum 17:5 et cecide- 17:5 And there fell ten portions to Manasseh, beside the

land of Gilead and Bashan, which were on the other side Jordan;

runt funiculi Manasse decem absque terra Galaad et Basan trans Iordanem 17:6 filiae enim Manasse possederunt he- 17:6 Because the daughters of Manasseh had an inheri-

tance among his sons: and the rest of Manasseh’s sons had the land of Gilead.

reditatem in medio filiorum eius terra autem Galaad cecidit in sortem filiorum Manasse qui reliqui erant 17:7 fuitque 17:7 And the coast of Manasseh was from Asher to Mich-

methah, that lieth before Shechem; and the border went along on the right hand unto the inhabitants of Entappuah.

Biblia Sacra Vulgata

terminus Manasse ab Aser Machmathath quae respicit Sychem et egreditur ad dextram iuxta habitatores fontis Taf-

17:8 Now Manasseh had the land of Tappuah: but Tap- fuae 17:8 etenim in sorte Manasse ceciderat terra Taffuae

puah on the border of Manasseh belonged to the children of Ephraim;

quae est iuxta terminos Manasse filiorum Ephraim 17:9 de-

17:9 And the coast descended unto the river Kanah, southward of the river: these cities of Ephraim are among the cities of Manasseh: the coast of Manasseh also was on the north side of the river, and the outgoings of it were at the sea:

scenditque terminus uallis Harundineti in meridiem torrentis ciuitatum Ephraim quae in medio sunt urbium Manasse terminus Manasse ab aquilone torrentis et exitus eius pergit ad

17:10 Southward it was Ephraim’s, and northward it was mare 17:10 ita ut ab austro sit possessio Ephraim et ab aqui-

Manasseh’s, and the sea is his border; and they met together in Asher on the north, and in Issachar on the east.

lone Manasse et utramque claudat mare et coniungantur sibi

17:11 And Manasseh had in Issachar and in Asher in tribu Aser ab aquilone et in tribu Isachar ab oriente 17:11

Bethshean and her towns, and Ibleam and her towns, and the inhabitants of Dor and her towns, and the inhabitants of Endor and her towns, and the inhabitants of Taanach and her towns, and the inhabitants of Megiddo and her towns, even three countries.

fuitque hereditas Manasse in Isachar et in Aser Bethsan et uiculi eius et Ieblaam cum uillulis suis et habitatores Dor cum oppidis suis habitatores quoque Hendor cum uillulis suis similiterque habitatores Thanach cum uillulis suis et habitatores Mageddo cum uiculis suis et tertia pars urbis Nofeth

17:12 Yet the children of Manasseh could not drive out the 17:12 nec potuerunt filii Manasse has subuertere ciuitates sed

inhabitants of those cities; but the Canaanites would dwell in that land.

coepit Chananeus habitare in terra ista 17:13 postquam au-

17:13 Yet it came to pass, when the children of Israel were waxen strong, that they put the Canaanites to tribute, but did not utterly drive them out.

tem conualuerunt filii Israhel subiecerunt Chananeos et fe-

17:14 And the children of Joseph spake unto Joshua, say- cerunt sibi tributarios nec interfecerunt eos 17:14 locutique

ing, Why hast thou given me but one lot and one portion to inherit, seeing I am a great people, forasmuch as the

LORD hath blessed me hitherto? sunt filii Ioseph ad Iosue atque dixerunt quare dedisti mihi

possessionem sortis et funiculi unius cum sim tantae mul-

17:15 And Joshua answered them, If thou be a great peo- titudinis et benedixerit mihi Dominus 17:15 ad quos Iosue

ple, then get thee up to the wood country, and cut down for thyself there in the land of the Perizzites and of the giants,

if mount Ephraim be too narrow for thee. ait si populus multus es ascende in siluam et succide tibi

spatia in terra Ferezei et Rafaim quia angusta est tibi pos-

17:16 And the children of Joseph said, The hill is not sessio montis Ephraim 17:16 cui responderunt filii Ioseph

enough for us: and all the Canaanites that dwell in the land of the valley have chariots of iron, both they who are of Bethshean and her towns, and they who are of the valley of Jezreel.

non poterimus ad montana conscendere cum ferreis curribus utantur Chananei qui habitant in terra campestri in qua sitae sunt Bethsan cum uiculis suis et Iezrahel mediam possidens

17:17 And Joshua spake unto the house of Joseph, even to uallem 17:17 dixitque Iosue ad domum Ioseph Ephraim et Ephraim and to Manasseh, saying, Thou art a great people, and hast great power: thou shalt not have one lot only:

Manasse populus multus es et magnae fortitudinis non ha-

17:18 But the mountain shall be thine; for it is a wood, bebis sortem unam 17:18 sed transibis ad montem et succi-

and thou shalt cut it down: and the outgoings of it shall be thine: for thou shalt drive out the Canaanites, though they

have iron chariots, and though they be strong. des tibi atque purgabis ad habitandum spatia et poteris ultra procedere cum subuerteris Chananeum quem dicis ferreos habere currus et esse fortissimum

18:1 congregatique sunt omnes filii Israhel in Silo ibique 18:1 And the whole congregation of the children of Israel

assembled together at Shiloh, and set up the tabernacle of the congregation there. And the land was subdued before

fixerunt tabernaculum testimonii et fuit eis terra subiecta them. 18:2 remanserant autem filiorum Israhel septem tribus quae 18:2 And there remained among the children of Israel

seven tribes, which had not yet received their inheritance.

necdum acceperant possessiones suas 18:3 ad quos Iosue 18:3 And Joshua said unto the children of Israel, How long are ye slack to go to possess the land, which the LORD God of your fathers hath given you?

ait usquequo marcetis ignauia et non intratis ad possidendam terram quam Dominus Deus patrum uestrorum dedit uobis 18:4 eligite de singulis tribubus ternos uiros ut mittam eos et 18:4 Give out from among you three men for each tribe:

and I will send them, and they shall rise, and go through the land, and describe it according to the inheritance of

pergant atque circumeant terram et describant eam iuxta nu- them; and they shall come again to me. merum uniuscuiusque multitudinis referantque ad me quod descripserint 18:5 diuidite uobis terram in septem partes Iu- 18:5 And they shall divide it into seven parts: Judah shall

abide in their coast on the south, and the house of Joseph shall abide in their coasts on the north.

das sit in terminis suis ab australi plaga et domus Ioseph ab

aquilone 18:6 mediam inter hos terram in septem partes de- 18:6 Ye shall therefore describe the land into seven parts, and bring the description hither to me, that I may cast lots for you here before the LORD our God.

scribite et huc uenietis ad me ut coram Domino Deo uestro mittam uobis hic sortem 18:7 quia non est inter uos pars 18:7 But the Levites have no part among you; for the

priesthood of the LORD is their inheritance: and Gad, and Reuben, and half the tribe of Manasseh, have received their inheritance beyond Jordan on the east, which Moses the servant of the LORD gave them.

Leuitarum sed sacerdotium Domini est eorum hereditas Gad autem et Ruben et dimidia tribus Manasse iam acceperant possessiones suas trans Iordanem ad orientalem plagam quas dedit eis Moses famulus Domini 18:8 cumque surrexissent 18:8 And the men arose, and went away: and Joshua

charged them that went to describe the land, saying, Go and walk through the land, and describe it, and come again to me, that I may here cast lots for you before the LORD in Shiloh.

uiri ut pergerent ad describendam terram praecepit eis Iosue dicens circuite terram et describite eam ac reuertimini ad me ut hic coram Domino Deo in Silo mittam uobis sortem 18:9 18:9 And the men went and passed through the land, and

described it by cities into seven parts in a book, and came again to Joshua to the host at Shiloh.

itaque perrexerunt et lustrantes eam in septem partes diuiserunt scribentes in uolumine reuersique sunt ad Iosue in castra Silo 18:10 qui misit sortes coram Domino in Silo diuisit- 18:10 And Joshua cast lots for them in Shiloh before the

LORD: and there Joshua divided the land unto the children of Israel according to their divisions.

que terram filiis Israhel in septem partes 18:11 et ascendit

18:11 And the lot of the tribe of the children of Benjamin came up according to their families: and the coast of their lot came forth between the children of Judah and the children of Joseph.

sors prima filiorum Beniamin per familias suas ut possiderent terram inter filios Iuda et filios Ioseph 18:12 fuitque 18:12 And their border on the north side was from Jordan;

and the border went up to the side of Jericho on the north side, and went up through the mountains westward; and

terminus eorum contra aquilonem ab Iordane pergens iuxta the goings out thereof were at the wilderness of Bethaven. latus Hiericho septentrionalis plagae et inde contra occidentem ad montana conscendens et perueniens in solitudinem Bethauen 18:13 atque pertransiens iuxta Luzam ad meri- 18:13 And the border went over from thence toward Luz,

to the side of Luz, which is Bethel, southward; and the border descended to Atarothadar, near the hill that lieth on

diem ipsa est Bethel descenditque in Atharothaddar in mon- the south side of the nether Bethhoron.

Biblia Sacra Vulgata

18:14 And the border was drawn thence, and compassed tem qui est ad meridiem Bethoron inferioris 18:14 et incli-

the corner of the sea southward, from the hill that lieth before Bethhoron southward; and the goings out thereof were at Kirjathbaal, which is Kirjathjearim, a city of the children of Judah: this was the west quarter.

natur circumiens contra mare a meridie montis qui respicit Bethoron contra africum suntque exitus eius in Cariathbaal quae uocatur et Cariathiarim urbem filiorum Iuda haec est

18:15 And the south quarter was from the end of Kirjath- plaga contra mare et occidentem 18:15 a meridie autem ex

jearim, and the border went out on the west, and went out to the well of waters of Nephtoah:

parte Cariathiarim egreditur terminus contra mare et perue-

18:16 And the border came down to the end of the moun- nit usque ad fontem aquarum Nepthoa 18:16 descenditque

tain that lieth before the valley of the son of Hinnom, and which is in the valley of the giants on the north, and descended to the valley of Hinnom, to the side of Jebusi on the south, and descended to Enrogel,

in partem montis qui respicit uallem filiorum Ennom et est contra septentrionalem plagam in extrema parte uallis Rafaim descenditque Gehennom id est uallis Ennom iuxta la-

18:17 And was drawn from the north, and went forth to tus Iebusei ad austrum et peruenit ad fontem Rogel 18:17

Enshemesh, and went forth toward Geliloth, which is over against the going up of Adummim, and descended to the

stone of Bohan the son of Reuben, transiens ad aquilonem et egrediens ad Aensemes id est fon-18:18 And passed along toward the side over against tem Solis 18:18 et pertransit usque ad tumulos qui sunt e

Arabah northward, and went down unto Arabah:

regione ascensus Adommim descenditque ad Abenboen id est lapidem Boen filii Ruben et pertransit ex latere aquilonis

18:19 And the border passed along to the side of Beth- ad campestria descenditque in planitiem 18:19 et praeterg-

hoglah northward: and the outgoings of the border were at the north bay of the salt sea at the south end of Jordan:

this was the south coast. reditur contra aquilonem Bethagla suntque exitus eius con-

tra linguam maris Salsissimi ab aquilone in fine Iordanis ad

18:20 And Jordan was the border of it on the east side. australem plagam 18:20 qui est terminus illius ab oriente

This was the inheritance of the children of Benjamin, by the coasts thereof round about, according to their families.

haec est possessio filiorum Beniamin per terminos suos in

18:21 Now the cities of the tribe of the children of Ben- circuitu et familias singulas 18:21 fueruntque ciuitates eius

jamin according to their families were Jericho, and Bethhoglah, and the valley of Keziz,

Hiericho et Bethagla et uallis Casis 18:22 Betharaba et Se-

18:22 And Betharabah, and Zemaraim, and Bethel,

18:23 And Avim, and Pharah, and Ophrah, maraim et Bethel 18:23 Auim et Affara et Ofra 18:24 uilla

18:24 And Chepharhaammonai, and Ophni, and Gaba; twelve cities with their villages:

Emona et Ofni et Gabee ciuitates duodecim et uillae earum

18:25 Gibeon, and Ramah, and Beeroth, 18:25 Gabaon et Rama et Beroth 18:26 et Mesfe Cafera et

18:26 And Mizpeh, and Chephirah, and Mozah,

18:27 And Rekem, and Irpeel, and Taralah, Ammosa 18:27 et Recem Iarafel et Tharala 18:28 et Sela

18:28 And Zelah, Eleph, and Jebusi, which is Jerusalem, Gibeath, and Kirjath; fourteen cities with their villages. This is the inheritance of the children of Benjamin according to their families.

Eleph et Iebus quae est Hierusalem Gabaath et Cariath ciuitates quattuordecim et uillae earum haec est possessio filiorum Beniamin iuxta familias suas

19:1 And the second lot came forth to Simeon, even for 19:1 et egressa est sors secunda filiorum Symeon per co-

the tribe of the children of Simeon according to their families: and their inheritance was within the inheritance of

the children of Judah. gnationes suas fuitque hereditas 19:2 eorum in medio pos-

19:2 And they had in their inheritance Beersheba, and Sheba, and Moladah,

sessionis filiorum Iuda Bersabee et Sabee et Molada 19:3

19:3 And Hazarshual, and Balah, and Azem,

19:4 And Eltolad, and Bethul, and Hormah, et Asersual Bala et Asem 19:4 et Heltholath Bethul Arma 19:5 et Seceleg et Bethmarchaboth Asersusa 19:6 et Beth- 19:5 And Ziklag, and Bethmarcaboth, and Hazarsusah,

19:6 And Bethlebaoth, and Sharuhen; thirteen cities and their villages:

lebaoth et Saroen ciuitates tredecim et uillae earum 19:7

19:7 Ain, Remmon, and Ether, and Ashan; four cities and their villages:

Ahin et Remmon et Athar et Asan ciuitates quattuor et uil-

lae earum 19:8 omnes uiculi per circuitum urbium istarum 19:8 And all the villages that were round about these cities to Baalathbeer, Ramath of the south. This is the inheritance of the tribe of the children of Simeon according to

usque ad Balaath Berrameth contra australem plagam haec their families. est hereditas filiorum Symeon iuxta cognationes suas 19:9 19:9 Out of the portion of the children of Judah was the

inheritance of the children of Simeon: for the part of the children of Judah was too much for them: therefore the children of Simeon had their inheritance within the inheritance of them.

in funiculo et possessione filiorum Iuda quia maior erat et idcirco possederunt filii Symeon in medio hereditatis eorum 19:10 cecidit quoque sors tertia filiorum Zabulon per cog- 19:10 And the third lot came up for the children of Ze-

bulun according to their families: and the border of their inheritance was unto Sarid:

nationes suas et factus est terminus possessionis eorum usque Sarith 19:11 ascenditque de mari et Medala ac peruenit 19:11 And their border went up toward the sea, and Mar-

alah, and reached to Dabbasheth, and reached to the river that is before Jokneam;

in Debbaseth usque ad torrentem qui est contra Iecennam 19:12 et reuertitur de Sarith contra orientem in fines Cese- 19:12 And turned from Sarid eastward toward the sunris-

ing unto the border of Chislothtabor, and then goeth out to Daberath, and goeth up to Japhia,

leththabor et egreditur ad Dabereth ascenditque contra Iafie 19:13 et inde pertransit ad orientalem plagam Getthefer Ett- 19:13 And from thence passeth on along on the east to

Gittahhepher, to Ittahkazin, and goeth out to Remmonmethoar to Neah;

hacasin et egreditur in Remmon Ampthar et Noa 19:14 et

19:14 And the border compasseth it on the north side to Hannathon: and the outgoings thereof are in the valley of Jiphthahel:

circuit ad aquilonem et Nathon suntque egressus eius uallis Iepthahel 19:15 et Catheth et Nehalal et Semron et Iedala 19:15 And Kattath, and Nahallal, and Shimron, and

Idalah, and Bethlehem: twelve cities with their villages.

et Bethleem ciuitates duodecim et uillae earum 19:16 haec 19:16 This is the inheritance of the children of Zebulun

according to their families, these cities with their villages.

est hereditas tribus filiorum Zabulon per cognationes suas urbes et uiculi earum 19:17 Isachar egressa est sors quarta 19:17 And the fourth lot came out to Issachar, for the chil-

dren of Issachar according to their families.

per cognationes suas 19:18 fuitque eius hereditas Hiezrahel 19:18 And their border was toward Jezreel, and Chesul-

loth, and Shunem,

et Chasaloth et Sunem 19:19 et Afaraim Seon et Anaarath 19:19 And Haphraim, and Shihon, and Anaharath, 19:20 et Rabbith et Cesion Abes 19:21 et Rameth et Engan- 19:20 And Rabbith, and Kishion, and Abez,

19:21 And Remeth, and Engannim, and Enhaddah, and Bethpazzez;

nim et Enadda et Bethfeses 19:22 et peruenit terminus us-

19:22 And the coast reacheth to Tabor, and Shahazimah, and Bethshemesh; and the outgoings of their border were

que Thabor et Seesima et Bethsemes eruntque exitus eius _{at Jordan: sixteen cities with their villages.} Iordanes ciuitates sedecim et uillae earum 19:23 haec est 19:23 This is the inheritance of the tribe of the children

of Issachar according to their families, the cities and their villages.

possessio filiorum Isachar per cognationes suas urbes et uiculi earum 19:24 cecidit sors quinta tribui filiorum Aser per 19:24 And the fifth lot came out for the tribe of the chil-

dren of Asher according to their families.

cognationes suas 19:25 fuitque terminus eorum Alchath et 19:25 And their border was Helkath, and Hali, and Beten,

and Achshaph,

Oali et Beten et Axab 19:26 Elmelech et Amaad et Mes- 19:26 And Alammelech, and Amad, and Misheal; and rea-

cheth to Carmel westward, and to Shihorlibnath;

sal et peruenit usque ad Carmelum maris et Siorlabanath 19:27 ac reuertitur contra orientem Bethdagon et pertransit 19:27 And turneth toward the sunrising to Bethdagon, and

reacheth to Zebulun, and to the valley of Jiphthahel toward the north side of Bethemek, and Neiel, and goeth out to Cabul on the left hand,

Biblia Sacra Vulgata

usque Zabulon et uallem Iepthahel contra aquilonem in Be-

19:28 And Hebron, and Rehob, and Hammon, and Kanah, themech et Neihel egrediturque ad leuam Chabul 19:28 et

even unto great Zidon;

Achran et Roob et Amon et Canae usque ad Sidonem mag-

19:29 And then the coast turneth to Ramah, and to the nam 19:29 reuertiturque in Orma usque ad ciuitatem muni-

strong city Tyre; and the coast turneth to Hosah; and the outgoings thereof are at the sea from the coast to Achzib:

tissimam Tyrum et usque Osa eruntque exitus eius in mare

19:30 Ummah also, and Aphek, and Rehob: twenty and de funiculo Acziba 19:30 et Amma et Afec et Roob ciui-

two cities with their villages.

19:31 This is the inheritance of the tribe of the children tates uiginti duae et uillae earum 19:31 haec est possessio

of Asher according to their families, these cities with their villages.

filiorum Aser per cognationes suas urbes et uiculi earum

19:32 The sixth lot came out to the children of Naphtali, 19:32 filiorum Nepthalim sexta pars cecidit per familias suas

even for the children of Naphtali according to their families.

19:33 et coepit terminus de Heleb et Helon in Sananim et

19:33 And their coast was from Heleph, from Allon to Zaanannim, and Adami, Nekeb, and Jabneel, unto Lakum; and the outgoings thereof were at Jordan:

Adami quae est Neceb et Iebnahel usque Lecum et egres-

19:34 And then the coast turneth westward to Aznothta- sus eorum usque ad Iordanem 19:34 reuertiturque terminus

bor, and goeth out from thence to Hukkok, and reacheth to Zebulun on the south side, and reacheth to Asher on the

west side, and to Judah upon Jordan toward the sunrising. contra occidentem in Aznoththabor atque inde egreditur in Vcoca et pertransit in Zabulon contra meridiem et in Aser contra occidentem et in Iuda ad Iordanem contra ortum so-

19:35 And the fenced cities are Ziddim, Zer, and Ham- lis 19:35 ciuitates munitissimae Aseddim Ser et Ammath et

math, Rakkath, and Chinnereth,

19:36 And Adamah, and Ramah, and Hazor, Recchath Chenereth 19:36 et Edema et Arama Asor 19:37

19:37 And Kedesh, and Edrei, and Enhazor, 19:38 And Iron, and Migdalel, Horem, and Bethanath, and et Cedes et Edrai Nasor 19:38 et Ieron et Magdalel Horem Bethshemesh; nineteen cities with their villages.

et Bethanath et Bethsemes ciuitates decem et nouem et uil-

19:39 This is the inheritance of the tribe of the children lae earum 19:39 haec est possessio tribus filiorum Nepthali

of Naphtali according to their families, the cities and their villages.

per cognationes suas urbes et uiculi earum 19:40 tribui fi-

19:40 And the seventh lot came out for the tribe of the children of Dan according to their families.

19:41 And the coast of their inheritance was Zorah, and liorum Dan per familias suas egressa est sors septima 19:41

Eshtaol, and Irshemesh,

et fuit terminus possessionis eius Saraa et Esthaol et Ahir-

19:42 And Shaalabbin, and Ajalon, and Jethlah, semes id est ciuitas Solis 19:42 Selebin et Ahialon et Iet-19:43 And Elon, and Thimnathah, and Ekron, hela 19:43 Helon et Themna et Acron 19:44 Helthecen et

19:44 And Eltekeh, and Gibbethon, and Baalath,

19:45 And Jehud, and Beneberak, and Gathrimmon, Gebthon et Baalath 19:45 Iud et Benebarach et Gethrem-

19:46 And Mejarkon, and Rakkon, with the border before mon 19:46 aquae Hiercon et Areccon cum termino qui re-Japho.

19:47 And the coast of the children of Dan went out too spicit Ioppen 19:47 et ipso fine concluditur ascenderuntque

little for them: therefore the children of Dan went up to fight against Leshem, and took it, and smote it with the edge of the sword, and possessed it, and dwelt therein, and called Leshem, Dan, after the name of Dan their father.

filii Dan et pugnauerunt contra Lesem ceperuntque eam et percusserunt in ore gladii ac possederunt et habitauerunt in ea uocantes nomen eius Lesemdan ex nomine Dan patris sui

19:48 This is the inheritance of the tribe of the children 19:48 haec est possessio tribus filiorum Dan per cognatio-

of Dan according to their families, these cities with their villages.

nes suas urbes et uiculi earum 19:49 cumque conplesset ter-

19:49 When they had made an end of dividing the land for inheritance by their coasts, the children of Israel gave an inheritance to Joshua the son of Nun among them:

ram sorte diuidere singulis per tribus suas dederunt filii Israhel possessionem Iosue filio Nun in medio sui 19:50 iuxta 19:50 According to the word of the LORD they gave

him the city which he asked, even Timnathserah in mount Ephraim: and he built the city, and dwelt therein.

praeceptum Domini urbem quam postulauit Thamnathseraa in monte Ephraim et aedificauit ciuitatem habitauitque in ea 19:51 hae sunt possessiones quas sorte diuiserunt Eleazar sa- 19:51 These are the inheritances, which Eleazar the priest,

and Joshua the son of Nun, and the heads of the fathers of the tribes of the children of Israel, divided for an inheritance by lot in Shiloh before the LORD, at the door of the tabernacle of the congregation. So they made an end of dividing the country.

cerdos et Iosue filius Nun et principes familiarum ac tribuum filiorum Israhel in Silo coram Domino ad ostium tabernaculi testimonii partitique sunt terram

20:1 et locutus est Dominus ad Iosue dicens loquere filiis 20:1 The LORD also spake unto Joshua, saying, Israhel et dic eis 20:2 separate urbes fugitiuorum de quibus 20:2 Speak to the children of Israel, saying, Appoint out

for you cities of refuge, whereof I spake unto you by the hand of Moses:

locutus sum ad uos per manum Mosi 20:3 ut confugiat ad

20:3 That the slayer that killeth any person unawares and unwittingly may flee thither: and they shall be your refuge from the avenger of blood.

eas quicumque animam percusserit nescius et possit euadere iram proximi qui ultor est sanguinis 20:4 cum ad unam ha- 20:4 And when he that doth flee unto one of those cities

shall stand at the entering of the gate of the city, and shall declare his cause in the ears of the elders of that city, they shall take him into the city unto them, and give him a place, that he may dwell among them.

rum confugerit ciuitatum stabitque ante portam ciuitatis et loquetur senioribus urbis illius ea quae se conprobent innocentem sicque suscipient eum et dabunt ei locum ad habitandum 20:5 cumque ultor sanguinis eum fuerit persecu- 20:5 And if the avenger of blood pursue after him, then

they shall not deliver the slayer up into his hand; because he smote his neighbour unwittingly, and hated him not be-

tus non tradent in manus eius quia ignorans percussit proxi- foretime. mum eius nec ante biduum triduumue eius probatur inimicus 20:6 et habitabit in ciuitate illa donec stet ante iudicium cau- 20:6 And he shall dwell in that city, until he stand be-

fore the congregation for judgment, and until the death of the high priest that shall be in those days: then shall the slayer return, and come unto his own city, and unto his own house, unto the city from whence he fled.

sam reddens facti sui et moriatur sacerdos magnus qui fuerit in illo tempore tunc reuertetur homicida et ingredietur ciuitatem et domum suam de qua fugerat 20:7 decreueruntque 20:7 And they appointed Kedesh in Galilee in mount

Naphtali, and Shechem in mount Ephraim, and Kirjatharba, which is Hebron, in the mountain of Judah.

Cedes in Galilea montis Nepthali et Sychem in monte Ephraim et Cariatharbe ipsa est Hebron in monte Iuda 20:8 et 20:8 And on the other side Jordan by Jericho eastward,

they assigned Bezer in the wilderness upon the plain out of the tribe of Reuben, and Ramoth in Gilead out of the tribe

trans Iordanem contra orientalem plagam Hiericho statue- of Gad, and Golan in Bashan out of the tribe of Manasseh. runt Bosor quae sita est in campestri solitudine de tribu Ruben et Ramoth in Galaad de tribu Gad et Gaulon in Basan de tribu Manasse 20:9 hae ciuitates constitutae sunt cunctis fi- 20:9 These were the cities appointed for all the children

of Israel, and for the stranger that sojourneth among them, that whosoever killeth any person at unawares might flee thither, and not die by the hand of the avenger of blood, until he stood before the congregation.

liis Israhel et aduenis qui habitant inter eos ut fugeret ad eas qui animam nescius percussisset et non moreretur in manu proximi effusum sanguinem uindicare cupientis donec staret ante populum expositurus causam suam

Biblia Sacra Vulgata

21:1 Then came near the heads of the fathers of the Levites 21:1 accesseruntque principes familiarum Leui ad Eleaunto Eleazar the priest, and unto Joshua the son of Nun, and unto the heads of the fathers of the tribes of the chil-

dren of Israel; zar sacerdotem et Iosue filium Nun et ad duces cognationum 21:2 And they spake unto them at Shiloh in the land of per singulas tribus filiorum Israhel 21:2 locutique sunt ad

Canaan, saying, The LORD commanded by the hand of Moses to give us cities to dwell in, with the suburbs thereof

for our cattle. eos in Silo terrae Chanaan atque dixerunt Dominus praece-

pit per manum Mosi ut darentur nobis urbes ad habitandum

21:3 And the children of Israel gave unto the Levites out et suburbana earum ad alenda iumenta 21:3 dederuntque of their inheritance, at the commandment of the LORD, these cities and their suburbs.

filii Israhel de possessionibus suis iuxta imperium Domini

21:4 And the lot came out for the families of the Ko- ciuitates et suburbana earum 21:4 egressaque est sors in

hathites: and the children of Aaron the priest, which were of the Levites, had by lot out of the tribe of Judah, and out of the tribe of Simeon, and out of the tribe of Benjamin, thirteen cities.

familiam Caath filiorum Aaron sacerdotis de tribubus Iuda

21:5 And the rest of the children of Kohath had by lot out et Symeon et Beniamin ciuitates tredecim 21:5 et reliquis

of the families of the tribe of Ephraim, and out of the tribe of Dan, and out of the half tribe of Manasseh, ten cities.

filiorum Caath id est Leuitis qui superflui erant de tribubus Ephraim et Dan et dimidia tribu Manasse ciuitates de-

21:6 And the children of Gershon had by lot out of the cem 21:6 porro filiis Gerson egressa est sors ut acciperent

families of the tribe of Issachar, and out of the tribe of Asher, and out of the tribe of Naphtali, and out of the half

tribe of Manasseh in Bashan, thirteen cities. de tribubus Isachar et Aser et Nepthalim dimidiaque tribu 21:7 The children of Merari by their families had out of Manasse in Basan ciuitates numero tredecim 21:7 et filiis

the tribe of Reuben, and out of the tribe of Gad, and out of the tribe of Zebulun, twelve cities.

Merari per cognationes suas de tribubus Ruben et Gad et

21:8 And the children of Israel gave by lot unto the Levites Zabulon urbes duodecim 21:8 dederuntque filii Israhel Lethese cities with their suburbs, as the LORD commanded by the hand of Moses.

uitis ciuitates et suburbana earum sicut praecepit Dominus

21:9 And they gave out of the tribe of the children of Ju- per manum Mosi singulis sorte tribuentes 21:9 de tribubus

dah, and out of the tribe of the children of Simeon, these cities which are here mentioned by name.

filiorum Iuda et Symeon dedit Iosue ciuitates quarum ista

21:10 Which the children of Aaron, being of the families sunt nomina 21:10 filiis Aaron per familias Caath leuitici

of the Kohathites, who were of the children of Levi, had: for theirs was the first lot.

generis prima enim sors illis egressa est 21:11 Cariatharbe

21:11 And they gave them the city of Arba the father of Anak, which city is Hebron, in the hill country of Judah, with the suburbs thereof round about it.

patris Enach quae uocatur Hebron in monte Iuda et subur-

21:12 But the fields of the city, and the villages thereof, bana eius per circuitum 21:12 agros uero et uillas eius de-

gave they to Caleb the son of Jephunneh for his possession.

derat Chaleb filio Iepphonne ad possidendum 21:13 dedit

21:13 Thus they gave to the children of Aaron the priest Hebron with her suburbs, to be a city of refuge for the slayer; and Libnah with her suburbs,

ergo filiis Aaron sacerdotis Hebron confugii ciuitatem ac su-

21:14 And Jattir with her suburbs, and Eshtemoa with her burbana eius et Lebnam cum suburbanis suis 21:14 et Iether

suburbs,

21:15 And Holon with her suburbs, and Debir with her et Isthimon 21:15 et Helon Dabir 21:16 et Ahin et Iethan

suburbs, 21:16 And Ain with her suburbs, and Juttah with her suburbs, and Bethshemesh with her suburbs; nine cities out of those two tribes.

et Bethsemes cum suburbanis suis ciuitates nouem de tribu-

21:17 And out of the tribe of Benjamin, Gibeon with her bus ut dictum est duabus 21:17 de tribu autem filiorum Be-

suburbs, Geba with her suburbs, 21:18 Anathoth with her suburbs, and Almon with her sub- niamin Gabaon et Gabee 21:18 et Anathoth et Almon cum urbs; four cities.

21:19 All the cities of the children of Aaron, the priests, suburbanis suis ciuitates quattuor 21:19 omnes simul ciuita-

were thirteen cities with their suburbs.

tes filiorum Aaron sacerdotis tredecim cum suburbanis suis 21:20 reliquis uero per familias filiorum Caath leuitici generis 21:20 And the families of the children of Kohath, the

Levites which remained of the children of Kohath, even they had the cities of their lot out of the tribe of Ephraim.

haec est data possessio 21:21 de tribu Ephraim urbs confu-

21:21 For they gave them Shechem with her suburbs in mount Ephraim, to be a city of refuge for the slayer; and Gezer with her suburbs,

gii Sychem cum suburbanis suis in monte Ephraim et Gazer 21:22 et Cebsain et Bethoron cum suburbanis suis ciuitates 21:22 And Kibzaim with her suburbs, and Bethhoron with

her suburbs; four cities.

quattuor 21:23 de tribu quoque Dan Elthece et Gebbethon 21:23 And out of the tribe of Dan, Eltekeh with her sub-

urbs, Gibbethon with her suburbs,

21:24 et Ahialon et Gethremmon cum suburbanis suis ciuita- 21:24 Aijalon with her suburbs, Gathrimmon with her sub-

urbs; four cities.

tes quattuor 21:25 porro de dimidia tribu Manasse Thanach 21:25 And out of the half tribe of Manasseh, Tanach with

her suburbs, and Gathrimmon with her suburbs; two cities.

et Gethremmon cum suburbanis suis ciuitates duae 21:26 21:26 All the cities were ten with their suburbs for the

families of the children of Kohath that remained.

omnes ciuitates decem et suburbana earum datae sunt filiis Caath inferioris gradus 21:27 filiis quoque Gerson le- 21:27 And unto the children of Gershon, of the families of

the Levites, out of the other half tribe of Manasseh they gave Golan in Bashan with her suburbs, to be a city of refuge for the slayer; and Beeshterah with her suburbs; two cities.

uitici generis dedit de dimidia tribu Manasse confugii ciuitatem Gaulon in Basan et Bosram cum suburbanis suis ciuitates duas 21:28 porro de tribu Isachar Cesion et Dabereth 21:28 And out of the tribe of Issachar, Kishon with her

suburbs, Dabareh with her suburbs,

21:29 et Iaramoth et Engannim cum suburbanis suis ciuitates 21:29 Jarmuth with her suburbs, Engannim with her sub-

urbs; four cities.

quattuor 21:30 de tribu autem Aser Masal et Abdon 21:31 21:30 And out of the tribe of Asher, Mishal with her sub-

urbs, Abdon with her suburbs,

21:31 Helkath with her suburbs, and Rehob with her suburbs; four cities.

et Elacoth et Roob cum suburbanis suis ciuitates quattuor 21:32 de tribu quoque Nepthali ciuitatem confugii Cedes in 21:32 And out of the tribe of Naphtali, Kedesh in Galilee

with her suburbs, to be a city of refuge for the slayer; and Hammothdor with her suburbs, and Kartan with her sub-

Galilea et Ammothdor et Charthan cum suburbanis suis ci- urbs; three cities. uitates tres 21:33 omnes urbes familiarum Gerson tredecim 21:33 All the cities of the Gershonites according to their

families were thirteen cities with their suburbs.

cum suburbanis suis 21:34 filiis autem Merari Leuitis infe- 21:34 And unto the families of the children of Merari, the

rest of the Levites, out of the tribe of Zebulun, Jokneam with her suburbs, and Kartah with her suburbs,

rioris gradus per familias suas data est de tribu Zabulon Iechenam et Chartha 21:35 et Damna et Nalol ciuitates quat- 21:35 Dimnah with her suburbs, Nahalal with her suburbs;

four cities.

tuor cum suburbanis suis 21:36 tribu quoque Ruben ciuita- 21:36 And out of the tribe of Reuben, Bezer with her sub-

urbs, and Jahazah with her suburbs, Kedemoth with her suburbs, and Mephaath with her suburbs; four cities.

tes confugii Bosor in solitudine et Cedson et Misor et Ocho ciuitates quattuor cum suburbanis 21:37 et de tribu Gad ci- 21:37 And out of the tribe of Gad, Ramoth in Gilead with

her suburbs, to be a city of refuge for the slayer; and Mahanaim with her suburbs, Heshbon with her suburbs, Jazer

uitates confugii Ramoth in Galaad et Manaim et Esebon et with her suburbs; four cities in all. Iazer ciuitates quattuor cum suburbanis suis 21:38 omnes 21:38 So all the cities for the children of Merari by their

families, which were remaining of the families of the Levites, were by their lot twelve cities.

urbes filiorum Merari per familias et cognationes suas duo-

decim 21:39 itaque uniuersae ciuitates Leuitarum in me- 21:39 All the cities of the Levites within the possession of the children of Israel were forty and eight cities with their suburbs.

dio possessionis filiorum Israhel fuerunt quadraginta octo 21:40 cum suburbanis suis singulae per familias distributae 21:40 These cities were every one with their suburbs round

about them: thus were all these cities.

21:41 deditque Dominus Israheli omnem terram quam tradi- 21:41 And the LORD gave unto Israel all the land which

he sware to give unto their fathers; and they possessed it, and dwelt therein.

turum se patribus eorum iurauerat et possederunt illam atque

Biblia Sacra Vulgata

21:42 And the LORD gave them rest round about, accord- habitauerunt in ea 21:42 dataque est ab eo pax in omnes per

ing to all that he sware unto their fathers: and there stood not a man of all their enemies before them; the LORD de-

livered all their enemies into their hand. circuitum nationes nullusque eis hostium resistere ausus est

21:43 There failed not ought of any good thing which the sed cuncti in eorum dicionem redacti sunt 21:43 ne unum LORD had spoken unto the house of Israel; all came to pass.

quidem uerbum quod illis praestaturum se esse promiserat irritum fuit sed rebus expleta sunt omnia

22:1 Then Joshua called the Reubenites, and the Gadites, 22:1 eodem tempore uocauit Iosue Rubenitas et Gadditas and the half tribe of Manasseh,

22:2 And said unto them, Ye have kept all that Moses the et dimidiam tribum Manasse 22:2 dixitque ad eos fecistis

servant of the LORD commanded you, and have obeyed my voice in all that I commanded you:

omnia quae uobis praecepit Moses famulus Domini mihi

22:3 Ye have not left your brethren these many days unto quoque in omnibus oboedistis 22:3 nec reliquistis fratres

this day, but have kept the charge of the commandment of the LORD your God.

uestros longo tempore usque in praesentem diem custodien-

22:4 And now the LORD your God hath given rest unto tes imperium Domini Dei uestri 22:4 quia igitur dedit Do-

your brethren, as he promised them: therefore now return ye, and get you unto your tents, and unto the land of your possession, which Moses the servant of the LORD gave you on the other side Jordan.

minus Deus uester fratribus uestris quietem ac pacem sicut pollicitus est reuertimini et ite in tabernacula uestra et in terram possessionis quam tradidit uobis Moses famulus

22:5 But take diligent heed to do the commandment and Domini trans Iordanem 22:5 ita dumtaxat ut custodiatis ad-

the law, which Moses the servant of the LORD charged you, to love the LORD your God, and to walk in all his ways, and to keep his commandments, and to cleave unto him, and to serve him with all your heart and with all your soul.

tente et opere conpleatis mandatum et legem quam praecepit uobis Moses seruus Domini ut diligatis Dominum Deum uestrum et ambuletis in omnibus uiis eius et obseruetis mandata illius adhereatisque ei ac seruiatis in omni corde et

22:6 So Joshua blessed them, and sent them away: and in omni anima uestra 22:6 benedixitque eis Iosue et dimi-

they went unto their tents.

22:7 Now to the one half of the tribe of Manasseh Moses sit eos qui reuersi sunt in tabernacula sua 22:7 tribui autem

had given possession in Bashan: but unto the other half thereof gave Joshua among their brethren on this side Jordan westward. And when Joshua sent them away also unto their tents, then he blessed them,

Manasse mediae possessionem Moses dederat in Basan et idcirco mediae quae superfuit dedit Iosue sortem inter ceteros fratres suos trans Iordanem ad occidentalem eius plagam cumque dimitteret eos in tabernacula sua et benedixisset il-

22:8 And he spake unto them, saying, Return with much lis 22:8 dixit ad eos in multa substantia atque diuitiis re-

riches unto your tents, and with very much cattle, with silver, and with gold, and with brass, and with iron, and with very much raiment: divide the spoil of your enemies with your brethren.

uertimini ad sedes uestras cum argento et auro aere ac ferro et ueste multiplici diuidite praedam hostium cum fratribus

22:9 And the children of Reuben and the children of Gad uestris 22:9 reuersique sunt et abierunt filii Ruben et filii

and the half tribe of Manasseh returned, and departed from the children of Israel out of Shiloh, which is in the land of Canaan, to go unto the country of Gilead, to the land of their possession, whereof they were possessed, according to the word of the LORD by the hand of Moses.

Gad et dimidia tribus Manasse a filiis Israhel de Silo quae sita est in Chanaan ut intrarent Galaad terram possessionis suae quam obtinuerant iuxta imperium Domini in manu

22:10 And when they came unto the borders of Jordan, Mosi 22:10 cumque uenissent ad tumulos Iordanis in terra

that are in the land of Canaan, the children of Reuben and the children of Gad and the half tribe of Manasseh built there an altar by Jordan, a great altar to see to.

Chanaan aedificauerunt iuxta Iordanem altare infinitae magnitudinis 22:11 quod cum audissent filii Israhel et ad eos 22:11 And the children of Israel heard say, Behold, the

children of Reuben and the children of Gad and the half tribe of Manasseh have built an altar over against the land of Canaan, in the borders of Jordan, at the passage of the children of Israel.

certi nuntii detulissent aedificasse filios Ruben et Gad et dimidiae tribus Manasse altare in terra Chanaan super Iordanis tumulos contra filios Israhel 22:12 conuenerunt omnes in 22:12 And when the children of Israel heard of it, the

whole congregation of the children of Israel gathered themselves together at Shiloh, to go up to war against

Silo ut ascenderent et dimicarent contra eos 22:13 et inte- them.

22:13 And the children of Israel sent unto the children of Reuben, and to the children of Gad, and to the half tribe of Manasseh, into the land of Gilead, Phinehas the son of Eleazar the priest,

rim miserunt ad illos in terram Galaad Finees filium Eleazar sacerdotem 22:14 et decem principes cum eo singulos de 22:14 And with him ten princes, of each chief house a

prince throughout all the tribes of Israel; and each one was an head of the house of their fathers among the thousands

tribubus singulis 22:15 qui uenerunt ad filios Ruben et Gad of Israel.

22:15 And they came unto the children of Reuben, and to the children of Gad, and to the half tribe of Manasseh, unto the land of Gilead, and they spake with them, saying,

et dimidiae tribus Manasse in terram Galaad dixeruntque ad eos 22:16 haec mandat omnis populus Domini quae est ista 22:16 Thus saith the whole congregation of the LORD,

What trespass is this that ye have committed against the God of Israel, to turn away this day from following the LORD, in that ye have builded you an altar, that ye might rebel this day against the LORD?

transgressio cur reliquistis Dominum Deum Israhel aedificantes altare sacrilegum et a cultu illius recedentes 22:17 22:17 Is the iniquity of Peor too little for us, from which

we are not cleansed until this day, although there was a plague in the congregation of the LORD,

an parum uobis est quod peccastis in Beelphegor et usque in praesentem diem macula huius sceleris in nobis permanet multique de populo corruerunt 22:18 et uos hodie reliquis- 22:18 But that ye must turn away this day from following

the LORD? and it will be, seeing ye rebel to day against the LORD, that to morrow he will be wroth with the whole

tis Dominum et cras in uniuersum Israhel eius ira desaeuiet congregation of Israel. 22:19 quod si putatis inmundam esse terram possessionis ue- 22:19 Notwithstanding, if the land of your possession be

unclean, then pass ye over unto the land of the possession of the LORD, wherein the LORD’s tabernacle dwelleth, and take possession among us: but rebel not against the LORD, nor rebel against us, in building you an altar beside the altar of the LORD our God.

strae transite ad terram in qua tabernaculum Domini est et habitate inter nos tantum ut a Domino et a nostro consortio non recedatis aedificato altari praeter altare Domini Dei uestri 22:20 nonne Achan filius Zare praeteriit mandatum 22:20 Did not Achan the son of Zerah commit a trespass in

the accursed thing, and wrath fell on all the congregation of Israel? and that man perished not alone in his iniquity.

Domini et super omnem populum Israhel ira eius incubuit et ille erat unus homo atque utinam solus perisset in scelere suo 22:21 responderuntque filii Ruben et Gad et dimidiae 22:21 Then the children of Reuben and the children of Gad

and the half tribe of Manasseh answered, and said unto the heads of the thousands of Israel,

tribus Manasse principibus legationis Israhel 22:22 fortis-

22:22 The LORD God of gods, the LORD God of gods, he knoweth, and Israel he shall know; if it be in rebellion, or if in transgression against the LORD, (save us not this day,)

simus Deus Dominus fortissimus Deus Dominus ipse nouit et Israhel simul intelleget si praeuaricationis animo hoc altare construximus non custodiat nos sed puniat in praesenti 22:23 et si ea mente fecimus ut holocausta et sacrificium et 22:23 That we have built us an altar to turn from following

the LORD, or if to offer thereon burnt offering or meat offering, or if to offer peace offerings thereon, let the LORD

pacificas uictimas super eo inponeremus ipse quaerat et iu- himself require it; dicet 22:24 et non ea magis cogitatione atque tractatu ut 22:24 And if we have not rather done it for fear of this

thing, saying, In time to come your children might speak unto our children, saying, What have ye to do with the

diceremus cras dicent filii uestri filiis nostris quid uobis et LORD God of Israel?

Biblia Sacra Vulgata

22:25 For the LORD hath made Jordan a border between Domino Deo Israhel 22:25 terminum posuit Dominus inter

us and you, ye children of Reuben and children of Gad; ye have no part in the LORD: so shall your children make

our children cease from fearing the LORD. nos et uos o filii Ruben et filii Gad Iordanem fluuium et idcirco partem non habetis in Domino et per hanc occasionem auertent filii uestri filios nostros a timore Domini putauimus

22:26 Therefore we said, Let us now prepare to build us itaque melius 22:26 et diximus extruamus nobis altare non

an altar, not for burnt offering, nor for sacrifice:

22:27 But that it may be a witness between us, and you, in holocausta neque ad uictimas offerendas 22:27 sed in te-

and our generations after us, that we might do the service of the LORD before him with our burnt offerings, and with our sacrifices, and with our peace offerings; that your children may not say to our children in time to come, Ye have no part in the LORD.

stimonium inter nos et uos et subolem nostram uestramque progeniem ut seruiamus Domino et iuris nostri sit offerre holocausta et uictimas et pacificas hostias et nequaquam dicant cras filii uestri filiis nostris non est uobis pars in Domino

22:28 Therefore said we, that it shall be, when they should 22:28 quod si uoluerint dicere respondebunt eis ecce altare

so say to us or to our generations in time to come, that we may say again, Behold the pattern of the altar of the LORD, which our fathers made, not for burnt offerings, nor for sacrifices; but it is a witness between us and you.

Domini quod fecerunt patres nostri non in holocausta neque

22:29 God forbid that we should rebel against the LORD, in sacrificium sed in testimonium uestrum ac nostrum 22:29

and turn this day from following the LORD, to build an altar for burnt offerings, for meat offerings, or for sacrifices, beside the altar of the LORD our God that is before his tabernacle.

absit a nobis hoc scelus ut recedamus a Domino et eius uestigia relinquamus extructo altari ad holocausta et sacrificia et uictimas offerendas praeter altare Domini Dei nostri quod

22:30 And when Phinehas the priest, and the princes of extructum est ante tabernaculum eius 22:30 quibus auditis

the congregation and heads of the thousands of Israel which were with him, heard the words that the children of Reuben and the children of Gad and the children of Manasseh spake, it pleased them.

Finees sacerdos et principes legationis Israhel qui erant cum eo placati sunt et uerba filiorum Ruben et Gad et dimidiae

22:31 And Phinehas the son of Eleazar the priest said unto tribus Manasse libentissime susceperunt 22:31 dixitque Fi-

the children of Reuben, and to the children of Gad, and to the children of Manasseh, This day we perceive that the LORD is among us, because ye have not committed this trespass against the LORD: now ye have delivered the children of Israel out of the hand of the LORD.

nees filius Eleazari sacerdos ad eos nunc scimus quod nobiscum sit Dominus quoniam alieni estis a praeuaricatione hac

22:32 And Phinehas the son of Eleazar the priest, and the et liberastis filios Israhel de manu Domini 22:32 reuersus-

princes, returned from the children of Reuben, and from the children of Gad, out of the land of Gilead, unto the land of Canaan, to the children of Israel, and brought them word again.

que est cum principibus a filiis Ruben et Gad de terra Galaad

22:33 And the thing pleased the children of Israel; and the finium Chanaan ad filios Israhel et rettulit eis 22:33 placuit-

children of Israel blessed God, and did not intend to go up against them in battle, to destroy the land wherein the

children of Reuben and Gad dwelt. que sermo cunctis audientibus et laudauerunt Deum filii Isra-

hel et nequaquam ultra dixerunt ut ascenderent contra eos at-

22:34 And the children of Reuben and the children of Gad que pugnarent et delerent terram possessionis eorum 22:34

called the altar Ed: for it shall be a witness between us that the LORD is God.

uocaueruntque filii Ruben et filii Gad altare quod extruxerant Testimonium nostrum quod Dominus ipse sit Deus

23:1 And it came to pass a long time after that the LORD 23:1 euoluto autem multo tempore postquam pacem Dohad given rest unto Israel from all their enemies round about, that Joshua waxed old and stricken in age.

minus dederat Israheli subiectis in gyro nationibus uniuersis

23:2 And Joshua called for all Israel, and for their elders, et Iosue iam longeuo et persenilis aetatis 23:2 uocauit Ioand for their heads, and for their judges, and for their officers, and said unto them, I am old and stricken in age:

sue omnem Israhelem maioresque natu et principes ac duces et magistros dixitque ad eos ego senui et progressioris aetatis sum 23:3 uosque cernitis omnia quae fecerit Dominus 23:3 And ye have seen all that the LORD your God hath

done unto all these nations because of you; for the LORD your God is he that hath fought for you.

Deus uester cunctis per circuitum nationibus quomodo pro uobis ipse pugnauerit 23:4 et nunc quia uobis sorte diui- 23:4 Behold, I have divided unto you by lot these nations

that remain, to be an inheritance for your tribes, from Jordan, with all the nations that I have cut off, even unto the

sit omnem terram ab orientali parte Iordanis usque ad mare great sea westward. Magnum multaeque adhuc supersunt nationes 23:5 Domi- 23:5 And the LORD your God, he shall expel them from

before you, and drive them from out of your sight; and ye shall possess their land, as the LORD your God hath

nus Deus uester disperdet eas et auferet a facie uestra et pos- promised unto you. sidebitis terram sicut uobis pollicitus est 23:6 tantum con- 23:6 Be ye therefore very courageous to keep and to do all

that is written in the book of the law of Moses, that ye turn not aside therefrom to the right hand or to the left;

fortamini et estote solliciti ut custodiatis cuncta quae scripta sunt in uolumine legis Mosi et non declinetis ab eis nec ad dextram nec ad sinistram 23:7 ne postquam intraueritis ad 23:7 That ye come not among these nations, these that re-

main among you; neither make mention of the name of their gods, nor cause to swear by them, neither serve them,

gentes quae inter uos futurae sunt iuretis in nomine deorum nor bow yourselves unto them: earum et seruiatis eis et adoretis illos 23:8 sed adhereatis 23:8 But cleave unto the LORD your God, as ye have done

unto this day.

Domino Deo uestro quod fecistis usque in diem hanc 23:9 23:9 For the LORD hath driven out from before you great

nations and strong: but as for you, no man hath been able to stand before you unto this day.

et tunc auferet Dominus in conspectu uestro gentes magnas et robustissimas et nullus uobis resistere poterit 23:10 unus 23:10 One man of you shall chase a thousand: for the

LORD your God, he it is that fighteth for you, as he hath promised you.

e uobis persequetur hostium mille uiros quia Dominus Deus uester pro uobis ipse pugnabit sicut pollicitus est 23:11 hoc 23:11 Take good heed therefore unto yourselves, that ye

love the LORD your God.

tantum diligentissime praecauete ut diligatis Dominum Deum uestrum 23:12 quod si uolueritis gentium harum quae 23:12 Else if ye do in any wise go back, and cleave

unto the remnant of these nations, even these that remain among you, and shall make marriages with them, and go

inter uos habitant erroribus adherere et cum eis miscere con- in unto them, and they to you: ubia atque amicitias copulare 23:13 iam nunc scitote quod 23:13 Know for a certainty that the LORD your God will

no more drive out any of these nations from before you; but they shall be snares and traps unto you, and scourges in your sides, and thorns in your eyes, until ye perish from off this good land which the LORD your God hath given you.

Dominus Deus uester non eas deleat ante faciem uestram sed sint uobis in foueam ac laqueum et offendiculum ex latere uestro et sudes in oculis uestris donec uos auferat atque disperdat de terra hac optima quam tradidit uobis 23:14 en 23:14 And, behold, this day I am going the way of all the

earth: and ye know in all your hearts and in all your souls, that not one thing hath failed of all the good things which the LORD your God spake concerning you; all are come to pass unto you, and not one thing hath failed thereof.

ego hodie ingrediar uiam uniuersae terrae et toto animo cognoscetis quod de omnibus uerbis quae se Dominus praestaturum nobis esse pollicitus est unum non praeterierit in cassum 23:15 sicut ergo impleuit opere quod promisit et pro- 23:15 Therefore it shall come to pass, that as all good

things are come upon you, which the LORD your God promised you; so shall the LORD bring upon you all evil things, until he have destroyed you from off this good land which the LORD your God hath given you.

spera cuncta uenerunt sic adducet super uos quicquid malorum comminatus est donec uos auferat atque disperdat de

Biblia Sacra Vulgata

23:16 When ye have transgressed the covenant of the terra hac optima quam tradidit uobis 23:16 eo quod praete-

LORD your God, which he commanded you, and have gone and served other gods, and bowed yourselves to them; then shall the anger of the LORD be kindled against you, and ye shall perish quickly from off the good land which he hath given unto you.

rieritis pactum Domini Dei uestri quod pepigit uobiscum et seruieritis diis alienis et adoraueritis eos cito atque uelociter consurget in uos furor Domini et auferemini de terra hac optima quam tradidit uobis

24:1 And Joshua gathered all the tribes of Israel to 24:1 congregauitque Iosue omnes tribus Israhel in Sychem

Shechem, and called for the elders of Israel, and for their heads, and for their judges, and for their officers; and they

presented themselves before God. et uocauit maiores natu ac principes et iudices et magistros 24:2 And Joshua said unto all the people, Thus saith the steteruntque in conspectu Domini 24:2 et ad populum sic

LORD God of Israel, Your fathers dwelt on the other side of the flood in old time, even Terah, the father of Abraham,

and the father of Nachor: and they served other gods. locutus est haec dicit Dominus Deus Israhel trans fluuium

habitauerunt patres uestri ab initio Thare pater Abraham et

24:3 And I took your father Abraham from the other side Nahor seruieruntque diis alienis 24:3 tuli ergo patrem ue-

of the flood, and led him throughout all the land of Canaan, and multiplied his seed, and gave him Isaac.

strum Abraham de Mesopotamiae finibus et adduxi eum in

24:4 And I gave unto Isaac Jacob and Esau: and I gave terram Chanaan multiplicauique semen eius 24:4 et dedi ei

unto Esau mount Seir, to possess it; but Jacob and his children went down into Egypt.

Isaac illique rursum dedi Iacob et Esau e quibus Esau dedi montem Seir ad possidendum Iacob uero et filii eius descen-

24:5 I sent Moses also and Aaron, and I plagued Egypt, derunt in Aegyptum 24:5 misique Mosen et Aaron et per-

according to that which I did among them: and afterward I brought you out.

cussi Aegyptum multis signis atque portentis 24:6 eduxi-

24:6 And I brought your fathers out of Egypt: and ye came unto the sea; and the Egyptians pursued after your fathers with chariots and horsemen unto the Red sea.

que uos et patres uestros de Aegypto et uenistis ad mare persecutique sunt Aegyptii patres uestros cum curribus et equi-

24:7 And when they cried unto the LORD, he put darkness tatu usque ad mare Rubrum 24:7 clamauerunt autem ad Do-

between you and the Egyptians, and brought the sea upon them, and covered them; and your eyes have seen what I have done in Egypt: and ye dwelt in the wilderness a long season.

minum filii Israhel qui posuit tenebras inter uos et Aegyptios et adduxit super eos mare et operuit illos uiderunt oculi uestri cuncta quae in Aegypto fecerim et habitastis in solitu-

24:8 And I brought you into the land of the Amorites, dine multo tempore 24:8 et introduxi uos ad terram Amor-

which dwelt on the other side Jordan; and they fought with you: and I gave them into your hand, that ye might possess

their land; and I destroyed them from before you. rei qui habitabat trans Iordanem cumque pugnarent contra

uos tradidi eos in manus uestras et possedistis terram eo-

24:9 Then Balak the son of Zippor, king of Moab, arose rum atque interfecistis illos 24:9 surrexit autem Balac filius

and warred against Israel, and sent and called Balaam the son of Beor to curse you:

Sepphor rex Moab et pugnauit contra Israhelem misitque et

24:10 But I would not hearken unto Balaam; therefore he uocauit Balaam filium Beor ut malediceret uobis 24:10 et

blessed you still: so I delivered you out of his hand.

ego nolui audire eum sed e contrario per illum benedixi uo-

24:11 And you went over Jordan, and came unto Jericho: bis et liberaui uos de manu eius 24:11 transistisque Iorda-

and the men of Jericho fought against you, the Amorites, and the Perizzites, and the Canaanites, and the Hittites, and the Girgashites, the Hivites, and the Jebusites; and I delivered them into your hand.

nem et uenistis ad Hiericho pugnaueruntque contra uos uiri ciuitatis eius Amorreus et Ferezeus et Chananeus et Hettheus et Gergeseus et Eueus et Iebuseus et tradidi illos in manus uestras 24:12 misique ante uos crabrones et eieci eos 24:12 And I sent the hornet before you, which drave them

out from before you, even the two kings of the Amorites; but not with thy sword, nor with thy bow.

de locis suis duos reges Amorreorum non in gladio et arcu tuo 24:13 dedique uobis terram in qua non laborastis et ur- 24:13 And I have given you a land for which ye did not

labour, and cities which ye built not, and ye dwell in them; of the vineyards and oliveyards which ye planted not do ye

bes quas non aedificastis ut habitaretis in eis uineas et oli- eat. ueta quae non plantastis 24:14 nunc ergo timete Dominum 24:14 Now therefore fear the LORD, and serve him in sin-

cerity and in truth: and put away the gods which your fathers served on the other side of the flood, and in Egypt;

et seruite ei perfecto corde atque uerissimo et auferte deos and serve ye the LORD. quibus seruierunt patres uestri in Mesopotamia et in Aegypto ac seruite Domino 24:15 sin autem malum uobis uidetur ut 24:15 And if it seem evil unto you to serve the LORD,

choose you this day whom ye will serve; whether the gods which your fathers served that were on the other side of the flood, or the gods of the Amorites, in whose land ye dwell: but as for me and my house, we will serve the LORD.

Domino seruiatis optio uobis datur eligite hodie quod placet cui potissimum seruire debeatis utrum diis quibus seruierunt patres uestri in Mesopotamia an diis Amorreorum in quorum terra habitatis ego autem et domus mea seruiemus Domino 24:16 responditque populus et ait absit a nobis ut relinqua- 24:16 And the people answered and said, God forbid that

we should forsake the LORD, to serve other gods;

mus Dominum et seruiamus diis alienis 24:17 Dominus De- 24:17 For the LORD our God, he it is that brought us up

and our fathers out of the land of Egypt, from the house of bondage, and which did those great signs in our sight, and preserved us in all the way wherein we went, and among all the people through whom we passed:

us noster ipse eduxit nos et patres nostros de terra Aegypti de domo seruitutis fecitque uidentibus nobis signa ingentia et custodiuit nos in omni uia per quam ambulauimus et in cunctis populis per quos transiuimus 24:18 et eiecit uniuersas 24:18 And the LORD drave out from before us all the peo-

ple, even the Amorites which dwelt in the land: therefore will we also serve the LORD; for he is our God.

gentes Amorreum habitatorem terrae quam nos intrauimus seruiemus igitur Domino quia ipse est Deus noster 24:19 24:19 And Joshua said unto the people, Ye cannot serve

the LORD: for he is an holy God; he is a jealous God; he will not forgive your transgressions nor your sins.

dixitque Iosue ad populum non poteritis seruire Domino Deus enim sanctus et fortis aemulator est nec ignoscet sceleribus uestris atque peccatis 24:20 si dimiseritis Dominum et 24:20 If ye forsake the LORD, and serve strange gods,

then he will turn and do you hurt, and consume you, after that he hath done you good.

seruieritis diis alienis conuertet se et adfliget uos atque subuertet postquam uobis praestiterit bona 24:21 dixitque po- 24:21 And the people said unto Joshua, Nay; but we will

serve the LORD.

pulus ad Iosue nequaquam ita ut loqueris erit sed Domino seruiemus 24:22 et Iosue ad populum testes inquit uos estis 24:22 And Joshua said unto the people, Ye are witnesses

against yourselves that ye have chosen you the LORD, to serve him. And they said, We are witnesses.

quia ipsi elegeritis uobis Dominum ut seruiatis ei responderuntque testes 24:23 nunc ergo ait auferte deos alienos de 24:23 Now therefore put away, said he, the strange gods

which are among you, and incline your heart unto the LORD God of Israel.

medio uestrum et inclinate corda uestra ad Dominum Deum Israhel 24:24 dixitque populus ad Iosue Domino Deo nostro 24:24 And the people said unto Joshua, The LORD our

God will we serve, and his voice will we obey.

seruiemus oboedientes praeceptis eius 24:25 percussit igi- 24:25 So Joshua made a covenant with the people that day,

and set them a statute and an ordinance in Shechem.

Biblia Sacra Vulgata

tur Iosue in die illo foedus et proposuit populo praecepta at-

24:26 And Joshua wrote these words in the book of the law que iudicia in Sychem 24:26 scripsitque omnia uerba haec

of God, and took a great stone, and set it up there under an oak, that was by the sanctuary of the LORD.

in uolumine legis Dei et tulit lapidem pergrandem posuitque

24:27 And Joshua said unto all the people, Behold, this eum subter quercum quae erat in sanctuario Domini 24:27

stone shall be a witness unto us; for it hath heard all the words of the LORD which he spake unto us: it shall be

therefore a witness unto you, lest ye deny your God. et dixit ad omnem populum en lapis iste erit uobis in testimonium quod audierit omnia uerba Domini quae locutus est uobis ne forte postea negare uelitis et mentiri Domino Deo

24:28 So Joshua let the people depart, every man unto his uestro 24:28 dimisitque populum singulos in possessionem

inheritance.

24:29 And it came to pass after these things, that Joshua suam 24:29 et post haec mortuus est Iosue filius Nun seruus

the son of Nun, the servant of the LORD, died, being an hundred and ten years old.

Domini centum decem annorum 24:30 sepelieruntque eum

24:30 And they buried him in the border of his inheritance in Timnathserah, which is in mount Ephraim, on the north side of the hill of Gaash.

in finibus possessionis suae in Thamnathsare quae sita est

24:31 And Israel served the LORD all the days of Joshua, in monte Ephraim a septentrionali parte montis Gaas 24:31

and all the days of the elders that overlived Joshua, and which had known all the works of the LORD, that he had

done for Israel. seruiuitque Israhel Domino cunctis diebus Iosue et seniorum

qui longo uixerunt tempore post Iosue et qui nouerant om-

24:32 And the bones of Joseph, which the children of Is- nia opera Domini quae fecerat in Israhel 24:32 ossa quoque

rael brought up out of Egypt, buried they in Shechem, in a parcel of ground which Jacob bought of the sons of Hamor the father of Shechem for an hundred pieces of silver: and it became the inheritance of the children of Joseph.

Ioseph quae tulerant filii Israhel de Aegypto sepelierunt in Sychem in parte agri quem emerat Iacob a filiis Emmor patris Sychem centum nouellis ouibus et fuit in possessione

24:33 And Eleazar the son of Aaron died; and they buried filiorum Ioseph 24:33 Eleazar quoque filius Aaron mortuus

him in a hill that pertained to Phinehas his son, which was given him in mount Ephraim.

est et sepelierunt eum in Gaab Finees filii eius quae data est ei in monte Ephraim

Liber Iudicum

1:1 post mortem Iosue consuluerunt filii Israhel Dominum 1:1 Now after the death of Joshua it came to pass, that

the children of Israel asked the LORD, saying, Who shall go up for us against the Canaanites first, to fight against

dicentes quis ascendet ante nos contra Chananeum et erit them? dux belli 1:2 dixitque Dominus Iudas ascendet ecce tradidi 1:2 And the LORD said, Judah shall go up: behold, I have

delivered the land into his hand.

terram in manus eius 1:3 et ait Iudas Symeoni fratri suo as- 1:3 And Judah said unto Simeon his brother, Come up with

me into my lot, that we may fight against the Canaanites; and I likewise will go with thee into thy lot. So Simeon

cende mecum in sorte mea et pugna contra Chananeum ut et went with him. ego pergam tecum in sorte tua et abiit cum eo Symeon 1:4 1:4 And Judah went up; and the LORD delivered the

Canaanites and the Perizzites into their hand: and they slew of them in Bezek ten thousand men.

ascenditque Iudas et tradidit Dominus Chananeum ac Ferezeum in manus eorum et percusserunt in Bezec decem milia uirorum 1:5 inueneruntque Adonibezec in Bezec et 1:5 And they found Adonibezek in Bezek: and they fought

against him, and they slew the Canaanites and the Perizzites.

pugnauerunt contra eum ac percusserunt Chananeum et Ferezeum 1:6 fugit autem Adonibezec quem secuti conpre- 1:6 But Adonibezek fled; and they pursued after him, and

caught him, and cut off his thumbs and his great toes.

henderunt caesis summitatibus manuum eius ac pedum 1:7 1:7 And Adonibezek said, Threescore and ten kings, hav-

ing their thumbs and their great toes cut off, gathered their meat under my table: as I have done, so God hath requited

dixitque Adonibezec septuaginta reges amputatis manuum me. And they brought him to Jerusalem, and there he died. ac pedum summitatibus colligebant sub mensa mea ciborum reliquias sicut feci ita reddidit mihi Deus adduxeruntque eum in Hierusalem et ibi mortuus est 1:8 obpugnantes ergo 1:8 Now the children of Judah had fought against

Jerusalem, and had taken it, and smitten it with the edge of the sword, and set the city on fire.

filii Iuda Hierusalem ceperunt eam et percusserunt in ore gladii tradentes cunctam incendio ciuitatem 1:9 et postea 1:9 And afterward the children of Judah went down to

fight against the Canaanites, that dwelt in the mountain, and in the south, and in the valley.

descendentes pugnauerunt contra Chananeum qui habitabat in montanis et ad meridiem et in campestribus 1:10 per- 1:10 And Judah went against the Canaanites that dwelt

in Hebron: (now the name of Hebron before was Kirjatharba:) and they slew Sheshai, and Ahiman, and Tal-

gensque Iudas contra Chananeum qui habitabat in Hebron mai. cui nomen fuit antiquitus Cariatharbe percussit Sisai et Ahiman et Tholmai 1:11 atque inde profectus abiit ad habitato- 1:11 And from thence he went against the inhabitants of

Debir: and the name of Debir before was Kirjathsepher:

res Dabir cuius nomen uetus erat Cariathsepher id est ciuitas Litterarum 1:12 dixitque Chaleb qui percusserit Cariathse- 1:12 And Caleb said, He that smiteth Kirjathsepher, and

taketh it, to him will I give Achsah my daughter to wife.

pher et uastauerit eam dabo ei Axam filiam meam uxorem 1:13 cumque cepisset eam Othonihel filius Cenez frater Cha- 1:13 And Othniel the son of Kenaz, Caleb’s younger

brother, took it: and he gave him Achsah his daughter to wife.

leb minor dedit ei filiam suam coniugem 1:14 quam per-

1:14 And it came to pass, when she came to him, that she moved him to ask of her father a field: and she lighted from off her ass; and Caleb said unto her, What wilt thou?

gentem in itinere monuit uir suus ut peteret a patre suo agrum quae cum suspirasset sedens asino dixit ei Chaleb quid habes 1:15 at illa respondit da mihi benedictionem quia ter- 1:15 And she said unto him, Give me a blessing: for thou

hast given me a south land; give me also springs of water. And Caleb gave her the upper springs and the nether

ram arentem dedisti mihi da et inriguam aquis dedit ergo ei springs. Chaleb inriguum superius et inriguum inferius 1:16 filii au- 1:16 And the children of the Kenite, Moses’ father in law,

went up out of the city of palm trees with the children of Judah into the wilderness of Judah, which lieth in the south of Arad; and they went and dwelt among the people.

Biblia Sacra Vulgata

tem Cinei cognati Mosi ascenderunt de ciuitate Palmarum cum filiis Iuda in desertum sortis eius quod est ad meri-

1:17 And Judah went with Simeon his brother, and they diem Arad et habitauerunt cum eo 1:17 abiit autem Iudas

slew the Canaanites that inhabited Zephath, and utterly destroyed it. And the name of the city was called Hormah.

cum Symeone fratre suo et percusserunt simul Chananeum qui habitabat in Sephath et interfecerunt eum uocatumque

1:18 Also Judah took Gaza with the coast thereof, and est nomen urbis Horma id est anathema 1:18 cepitque Iu-

Askelon with the coast thereof, and Ekron with the coast thereof.

das Gazam cum finibus suis et Ascalonem atque Accaron

1:19 And the LORD was with Judah; and he drave out cum terminis suis 1:19 fuitque Dominus cum Iuda et mon-

the inhabitants of the mountain; but could not drive out the inhabitants of the valley, because they had chariots of

iron. tana possedit nec potuit delere habitatores uallis quia falcatis 1:20 And they gave Hebron unto Caleb, as Moses said: curribus abundabant 1:20 dederuntque Chaleb Hebron sic-

and he expelled thence the three sons of Anak.

1:21 And the children of Benjamin did not drive out the ut dixerat Moses qui deleuit ex ea tres filios Enach 1:21 Ie-

Jebusites that inhabited Jerusalem; but the Jebusites dwell with the children of Benjamin in Jerusalem unto this day.

buseum autem habitatorem Hierusalem non deleuerunt filii Beniamin habitauitque Iebuseus cum filiis Beniamin in Hie-

1:22 And the house of Joseph, they also went up against rusalem usque in praesentem diem 1:22 domus quoque Io-

Bethel: and the LORD was with them.

1:23 And the house of Joseph sent to descry Bethel. (Now seph ascendit in Bethel fuitque Dominus cum eis 1:23 nam

the name of the city before was Luz.)

1:24 And the spies saw a man come forth out of the city, cum obsiderent urbem quae prius Luza uocabatur 1:24 ui-

and they said unto him, Shew us, we pray thee, the entrance into the city, and we will shew thee mercy.

derunt hominem egredientem de ciuitate dixeruntque ad eum ostende nobis introitum ciuitatis et faciemus tecum miseri-

1:25 And when he shewed them the entrance into the city, cordiam 1:25 qui cum ostendisset eis percusserunt urbem

they smote the city with the edge of the sword; but they let go the man and all his family.

in ore gladii hominem autem illum et omnem cognationem

1:26 And the man went into the land of the Hittites, and eius dimiserunt 1:26 qui dimissus abiit in terram Etthim et

built a city, and called the name thereof Luz: which is the name thereof unto this day.

aedificauit ibi ciuitatem uocauitque eam Luzam quae ita ap-

1:27 Neither did Manasseh drive out the inhabitants of pellatur usque in praesentem diem 1:27 Manasses quoque

Bethshean and her towns, nor Taanach and her towns, nor the inhabitants of Dor and her towns, nor the inhabitants of Ibleam and her towns, nor the inhabitants of Megiddo and her towns: but the Canaanites would dwell in that land.

non deleuit Bethsan et Thanach cum uiculis suis et habitatores Dor et Ieblaam et Mageddo cum uiculis suis coepitque

1:28 And it came to pass, when Israel was strong, that they Chananeus habitare cum eis 1:28 postquam autem conforput the Canaanites to tribute, and did not utterly drive them out.

tatus est Israhel fecit eos tributarios et delere noluit 1:29

1:29 Neither did Ephraim drive out the Canaanites that dwelt in Gezer; but the Canaanites dwelt in Gezer among them.

Ephraim etiam non interfecit Chananeum qui habitabat in

1:30 Neither did Zebulun drive out the inhabitants of Gazer sed habitauit cum eo 1:30 Zabulon non deleuit habi-

Kitron, nor the inhabitants of Nahalol; but the Canaanites dwelt among them, and became tributaries.

tatores Cetron et Naalon sed habitauit Chananeus in medio

1:31 Neither did Asher drive out the inhabitants of Accho, eius factusque est ei tributarius 1:31 Aser quoque non denor the inhabitants of Zidon, nor of Ahlab, nor of Achzib, nor of Helbah, nor of Aphik, nor of Rehob:

leuit habitatores Achcho et Sidonis Alab et Achazib et Alba

1:32 But the Asherites dwelt among the Canaanites, the et Afec et Roob 1:32 habitauitque in medio Chananei hainhabitants of the land: for they did not drive them out.

Liber Iudicum

bitatoris illius terrae nec interfecit eum 1:33 Nepthali non 1:33 Neither did Naphtali drive out the inhabitants of

Bethshemesh, nor the inhabitants of Bethanath; but he dwelt among the Canaanites, the inhabitants of the land: nevertheless the inhabitants of Bethshemesh and of Bethanath became tributaries unto them.

deleuit habitatores Bethsemes et Bethanath et habitauit inter Chananeum habitatorem terrae fueruntque ei Bethsemitae et Bethanitae tributarii 1:34 artauitque Amorreus filios Dan in 1:34 And the Amorites forced the children of Dan into the

mountain: for they would not suffer them to come down to the valley:

monte nec dedit eis locum ut ad planiora descenderent 1:35

1:35 But the Amorites would dwell in mount Heres in Aijalon, and in Shaalbim: yet the hand of the house of Joseph prevailed, so that they became tributaries.

habitauitque in monte Hares quod interpretatur testaceo in Ahilon et Salabim et adgrauata est manus domus Ioseph factusque est ei tributarius 1:36 fuit autem terminus Amorrei 1:36 And the coast of the Amorites was from the going up

to Akrabbim, from the rock, and upward.

ab ascensu Scorpionis Petra et superiora loca

2:1 ascenditque angelus Domini de Galgal ad locum Flen- 2:1 And an angel of the LORD came up from Gilgal to

Bochim, and said, I made you to go up out of Egypt, and have brought you unto the land which I sware unto your fathers; and I said, I will never break my covenant with you.

tium et ait eduxi uos de Aegypto et introduxi in terram pro qua iuraui patribus uestris et pollicitus sum ut non face-

rem irritum pactum meum uobiscum in sempiternum 2:2 2:2 And ye shall make no league with the inhabitants of this land; ye shall throw down their altars: but ye have not obeyed my voice: why have ye done this?

ita dumtaxat ut non feriretis foedus cum habitatoribus terrae huius et aras eorum subuerteretis et noluistis audire uo-

cem meam cur hoc fecistis 2:3 quam ob rem nolui delere 2:3 Wherefore I also said, I will not drive them out from before you; but they shall be as thorns in your sides, and their gods shall be a snare unto you.

eos a facie uestra ut habeatis hostes et dii eorum sint uobis in ruinam 2:4 cumque loqueretur angelus Domini uerba 2:4 And it came to pass, when the angel of the LORD

spake these words unto all the children of Israel, that the people lifted up their voice, and wept.

haec ad omnes filios Israhel eleuauerunt uocem suam et fleuerunt 2:5 et uocatum est nomen loci illius Flentium siue 2:5 And they called the name of that place Bochim: and

they sacrificed there unto the LORD.

Lacrimarum immolaueruntque ibi hostias Domino 2:6 di- 2:6 And when Joshua had let the people go, the children

of Israel went every man unto his inheritance to possess the land.

misit ergo Iosue populum et abierunt filii Israhel unusquisque in possessionem suam ut obtinerent eam 2:7 seruie- 2:7 And the people served the LORD all the days of

Joshua, and all the days of the elders that outlived Joshua, who had seen all the great works of the LORD, that he did

runtque Domino cunctis diebus eius et seniorum qui longo for Israel. post eum uixerunt tempore et nouerant omnia opera Domini quae fecerat cum Israhel 2:8 mortuus est autem Iosue fi- 2:8 And Joshua the son of Nun, the servant of the LORD,

died, being an hundred and ten years old.

lius Nun famulus Domini centum et decem annorum 2:9 et 2:9 And they buried him in the border of his inheritance in Timnathheres, in the mount of Ephraim, on the north side of the hill Gaash.

sepelierunt eum in finibus possessionis suae in Thamnathsare in monte Ephraim a septentrionali plaga montis Gaas 2:10 omnisque illa generatio congregata est ad patres suos 2:10 And also all that generation were gathered unto their

fathers: and there arose another generation after them, which knew not the LORD, nor yet the works which he

et surrexerunt alii qui non nouerant Dominum et opera quae had done for Israel. fecerat cum Israhel 2:11 feceruntque filii Israhel malum in 2:11 And the children of Israel did evil in the sight of the

LORD, and served Baalim:

conspectu Domini et seruierunt Baalim 2:12 ac dimiserunt 2:12 And they forsook the LORD God of their fathers,

which brought them out of the land of Egypt, and followed other gods, of the gods of the people that were round about them, and bowed themselves unto them, and provoked the LORD to anger.

Biblia Sacra Vulgata

Dominum Deum patrum suorum qui eduxerat eos de terra Aegypti et secuti sunt deos alienos deos quoque populorum qui habitabant in circuitu eorum et adorauerunt eos et ad ira-

2:13 And they forsook the LORD, and served Baal and cundiam concitauerunt Dominum 2:13 dimittentes eum et

Ashtaroth.

2:14 And the anger of the LORD was hot against Is- seruientes Baal et Astharoth 2:14 iratusque Dominus con-

rael, and he delivered them into the hands of spoilers that spoiled them, and he sold them into the hands of their enemies round about, so that they could not any longer stand before their enemies.

tra Israhel tradidit eos in manibus diripientium qui ceperunt eos et uendiderunt hostibus qui habitabant per gyrum nec po-

2:15 Whithersoever they went out, the hand of the LORD tuerunt resistere aduersariis suis 2:15 sed quocumque per-

was against them for evil, as the LORD had said, and as the LORD had sworn unto them: and they were greatly

distressed. gere uoluissent manus Domini erat super eos sicut locutus 2:16 Nevertheless the LORD raised up judges, which de- est et iurauit eis et uehementer adflicti sunt 2:16 suscitauit-

livered them out of the hand of those that spoiled them.

que Dominus iudices qui liberarent eos de uastantium ma-

2:17 And yet they would not hearken unto their judges, but nibus sed nec illos audire uoluerunt 2:17 fornicantes cum

they went a whoring after other gods, and bowed themselves unto them: they turned quickly out of the way which their fathers walked in, obeying the commandments of the LORD; but they did not so.

diis alienis et adorantes eos cito deseruerunt uiam per quam ingressi fuerant patres eorum et audientes mandata Domini

2:18 And when the LORD raised them up judges, then omnia fecere contraria 2:18 cumque Dominus iudices su-

the LORD was with the judge, and delivered them out of the hand of their enemies all the days of the judge: for it repented the LORD because of their groanings by reason of them that oppressed them and vexed them.

scitaret in diebus eorum flectebatur misericordia et audiebat adflictorum gemitus et liberabat eos de caede uastantium

2:19 And it came to pass, when the judge was dead, that 2:19 postquam autem mortuus esset iudex reuertebantur et

they returned, and corrupted themselves more than their fathers, in following other gods to serve them, and to bow down unto them; they ceased not from their own doings, nor from their stubborn way.

multo maiora faciebant quam fecerant patres sui sequentes deos alienos et seruientes eis et adorantes illos non dimiserunt adinuentiones suas et uiam durissimam per quam ambu-

2:20 And the anger of the LORD was hot against Israel; lare consueuerant 2:20 iratusque est furor Domini in Israhel

and he said, Because that this people hath transgressed my covenant which I commanded their fathers, and have not

hearkened unto my voice; et ait quia irritum fecit gens ista pactum meum quod pepige-

ram cum patribus eorum et uocem meam audire contempsit

2:21 I also will not henceforth drive out any from before 2:21 et ego non delebo gentes quas dimisit Iosue et mortuus

them of the nations which Joshua left when he died:

2:22 That through them I may prove Israel, whether they est 2:22 ut in ipsis experiar Israhel utrum custodiant uiam

will keep the way of the LORD to walk therein, as their fathers did keep it, or not.

Domini et ambulent in ea sicut custodierunt patres eorum an

2:23 Therefore the LORD left those nations, without driv- non 2:23 dimisit ergo Dominus omnes has nationes et cito ing them out hastily; neither delivered he them into the hand of Joshua.

subuertere noluit nec tradidit in manibus Iosue

3:1 Now these are the nations which the LORD left, to 3:1 hae sunt gentes quas Dominus dereliquit ut erudiret in

prove Israel by them, even as many of Israel as had not known all the wars of Canaan;

eis Israhelem et omnes qui non nouerant bella Chananeorum

3:2 Only that the generations of the children of Israel 3:2 et postea discerent filii eorum certare cum hostibus et ha-

might know, to teach them war, at the least such as before knew nothing thereof;

bere consuetudinem proeliandi 3:3 quinque satrapas Philis-

3:3 Namely, five lords of the Philistines, and all the Canaanites, and the Sidonians, and the Hivites that dwelt in mount Lebanon, from mount Baalhermon unto the entering in of Hamath.

Liber Iudicum

thinorum omnemque Chananeum et Sidonium atque Eueum qui habitabat in monte Libano de monte Baalhermon usque ad introitum Emath 3:4 dimisitque eos ut in ipsis experi- 3:4 And they were to prove Israel by them, to know

whether they would hearken unto the commandments of the LORD, which he commanded their fathers by the hand

retur Israhelem utrum audiret mandata Domini quae prae- of Moses.

ceperat patribus eorum per manum Mosi an non 3:5 ita- 3:5 And the children of Israel dwelt among the Canaanites, Hittites, and Amorites, and Perizzites, and Hivites, and Jebusites:

que filii Israhel habitauerunt in medio Chananei et Hetthei et Amorrei et Ferezei et Euei et Iebusei 3:6 et duxerunt uxo- 3:6 And they took their daughters to be their wives, and

gave their daughters to their sons, and served their gods.

res filias eorum ipsique filias suas eorum filiis tradiderunt et seruierunt diis eorum 3:7 feceruntque malum in conspectu 3:7 And the children of Israel did evil in the sight of

the LORD, and forgat the LORD their God, and served Baalim and the groves.

Domini et obliti sunt Dei sui seruientes Baalim et Astharoth 3:8 iratusque Dominus contra Israhel tradidit eos in 3:8 Therefore the anger of the LORD was hot against Is-

rael, and he sold them into the hand of Chushanrishathaim king of Mesopotamia: and the children of Israel served

manus Chusanrasathaim regis Mesopotamiae seruieruntque Chushanrishathaim eight years. ei octo annis 3:9 et clamauerunt ad Dominum qui suscita- 3:9 And when the children of Israel cried unto the LORD,

the LORD raised up a deliverer to the children of Israel, who delivered them, even Othniel the son of Kenaz,

uit eis saluatorem et liberauit eos Othonihel uidelicet filium Caleb’s younger brother. Cenez fratrem Chaleb minorem 3:10 fuitque in eo spiritus 3:10 And the Spirit of the LORD came upon him, and he

judged Israel, and went out to war: and the LORD delivered Chushanrishathaim king of Mesopotamia into his

Domini et iudicauit Israhel egressusque est ad pugnam et hand; and his hand prevailed against Chushanrishathaim.

tradidit Dominus in manu eius Chusanrasathaim regem Syriae et oppressit eum 3:11 quieuitque terra quadraginta an- 3:11 And the land had rest forty years. And Othniel the

son of Kenaz died.

nis et mortuus est Othonihel filius Cenez 3:12 addiderunt 3:12 And the children of Israel did evil again in the sight

of the LORD: and the LORD strengthened Eglon the king of Moab against Israel, because they had done evil in the

autem filii Israhel facere malum in conspectu Domini qui sight of the LORD. confortauit aduersum eos Eglon regem Moab quia fecerunt malum in conspectu eius 3:13 et copulauit ei filios Ammon 3:13 And he gathered unto him the children of Ammon

and Amalek, and went and smote Israel, and possessed the city of palm trees.

et Amalech abiitque et percussit Israhel atque possedit urbem Palmarum 3:14 seruieruntque filii Israhel Eglon regi 3:14 So the children of Israel served Eglon the king of

Moab eighteen years.

Moab decem et octo annis 3:15 et postea clamauerunt ad 3:15 But when the children of Israel cried unto the LORD, the LORD raised them up a deliverer, Ehud the son of Gera, a Benjamite, a man lefthanded: and by him the chil-

Dominum qui suscitauit eis saluatorem uocabulo Ahoth fi- dren of Israel sent a present unto Eglon the king of Moab. lium Gera filii Iemini qui utraque manu utebatur pro dextera miseruntque filii Israhel per illum munera Eglon regi Moab 3:16 qui fecit sibi gladium ancipitem habentem in medio ca- 3:16 But Ehud made him a dagger which had two edges,

of a cubit length; and he did gird it under his raiment upon his right thigh.

pulum longitudinis palmae manus et accinctus est eo subter sagum in dextro femore 3:17 obtulitque munera Eglon regi 3:17 And he brought the present unto Eglon king of Moab:

and Eglon was a very fat man.

Moab erat autem Eglon crassus nimis 3:18 cumque obtu- 3:18 And when he had made an end to offer the present,

he sent away the people that bare the present.

lisset ei munera prosecutus est socios qui cum eo uenerant

Biblia Sacra Vulgata

3:19 But he himself turned again from the quarries that 3:19 et reuersus de Galgalis ubi erant idola dixit ad regem

were by Gilgal, and said, I have a secret errand unto thee, O king: who said, Keep silence. And all that stood by him

went out from him. uerbum secretum habeo ad te o rex et ille imperauit silentium 3:20 And Ehud came unto him; and he was sitting in a egressisque omnibus qui circa eum erant 3:20 ingressus est

summer parlour, which he had for himself alone. And Ehud said, I have a message from God unto thee. And

he arose out of his seat. Ahoth ad eum sedebat autem in aestiuo cenaculo solus dixit-

que uerbum Dei habeo ad te qui statim surrexit de throno

3:21 And Ehud put forth his left hand, and took the dagger 3:21 extenditque Ahoth manum sinistram et tulit sicam de de-

from his right thigh, and thrust it into his belly: 3:22 And the haft also went in after the blade; and the fat xtro femore suo infixitque eam in uentre eius 3:22 tam uaclosed upon the blade, so that he could not draw the dagger out of his belly; and the dirt came out.

lide ut capulus ferrum sequeretur in uulnere ac pinguissimo adipe stringeretur nec eduxit gladium sed ita ut percusserat reliquit in corpore statimque per secreta naturae alui stercora

3:23 Then Ehud went forth through the porch, and shut the proruperunt 3:23 Ahoth autem clausis diligentissime ostiis

doors of the parlour upon him, and locked them.

3:24 When he was gone out, his servants came; and cenaculi et obfirmatis sera 3:24 per posticam egressus est

when they saw that, behold, the doors of the parlour were locked, they said, Surely he covereth his feet in his sum-

mer chamber. seruique regis ingressi uiderunt clausas fores cenaculi atque 3:25 And they tarried till they were ashamed: and, behold, dixerunt forsitan purgat aluum in aestiuo cubiculo 3:25 ex-

he opened not the doors of the parlour; therefore they took a key, and opened them: and, behold, their lord was fallen

down dead on the earth. pectantesque diu donec erubescerent et uidentes quod nullus

aperiret tulerunt clauem et aperientes inuenerunt dominum

3:26 And Ehud escaped while they tarried, and passed be- suum iacentem in terra mortuum 3:26 Ahoth autem dum yond the quarries, and escaped unto Seirath.

illi turbarentur effugit et pertransiit locum Idolorum unde

3:27 And it came to pass, when he was come, that he blew reuersus fuerat uenitque in Seirath 3:27 et statim insonuit

a trumpet in the mountain of Ephraim, and the children of Israel went down with him from the mount, and he before

them. bucina in monte Ephraim descenderuntque cum eo filii Isra-3:28 And he said unto them, Follow after me: for the hel ipso in fronte gradiente 3:28 qui dixit ad eos sequimini

LORD hath delivered your enemies the Moabites into your hand. And they went down after him, and took the fords of

Jordan toward Moab, and suffered not a man to pass over. me tradidit enim Dominus inimicos nostros Moabitas in manus nostras descenderuntque post eum et occupauerunt uada Iordanis quae transmittunt in Moab et non dimiserunt trans-

3:29 And they slew of Moab at that time about ten thou- ire quemquam 3:29 sed percusserunt Moabitas in tempore

sand men, all lusty, and all men of valour; and there escaped not a man.

illo circiter decem milia omnes robustos et fortes uiros nul-

3:30 So Moab was subdued that day under the hand of lus eorum euadere potuit 3:30 humiliatusque est Moab die

Israel. And the land had rest fourscore years.

3:31 And after him was Shamgar the son of Anath, which illo sub manu Israhel et quieuit terra octoginta annis 3:31

slew of the Philistines six hundred men with an ox goad: and he also delivered Israel.

post hunc fuit Samgar filius Anath qui percussit de Philisthim sescentos uiros uomere et ipse quoque defendit Israhel

4:1 And the children of Israel again did evil in the sight of 4:1 addideruntque filii Israhel facere malum in conspectu the LORD, when Ehud was dead.

4:2 And the LORD sold them into the hand of Jabin king Domini post mortem Ahoth 4:2 et tradidit illos Dominus

of Canaan, that reigned in Hazor; the captain of whose host was Sisera, which dwelt in Harosheth of the Gentiles.

in manu Iabin regis Chanaan qui regnauit in Asor habuitque

Liber Iudicum

ducem exercitus sui nomine Sisaram ipse autem habitabat in Aroseth gentium 4:3 clamaueruntque filii Israhel ad Domi- 4:3 And the children of Israel cried unto the LORD: for

he had nine hundred chariots of iron; and twenty years he mightily oppressed the children of Israel.

num nongentos enim habebat falcatos currus et per uiginti annos uehementer oppresserat eos 4:4 erat autem Debbora 4:4 And Deborah, a prophetess, the wife of Lapidoth, she

judged Israel at that time.

prophetis uxor Lapidoth quae iudicabat populum in illo tempore 4:5 et sedebat sub palma quae nomine illius uocabatur 4:5 And she dwelt under the palm tree of Deborah between

Ramah and Bethel in mount Ephraim: and the children of Israel came up to her for judgment.

inter Rama et Bethel in monte Ephraim ascendebantque ad eam filii Israhel in omne iudicium 4:6 quae misit et uocauit 4:6 And she sent and called Barak the son of Abinoam out

of Kedeshnaphtali, and said unto him, Hath not the LORD God of Israel commanded, saying, Go and draw toward mount Tabor, and take with thee ten thousand men of the children of Naphtali and of the children of Zebulun?

Barac filium Abinoem de Cedes Nepthalim dixitque ad eum praecepit tibi Dominus Deus Israhel uade et duc exercitum in montem Thabor tollesque tecum decem milia pugnatorum de filiis Nepthalim et de filiis Zabulon 4:7 ego autem 4:7 And I will draw unto thee to the river Kishon Sisera,

the captain of Jabin’s army, with his chariots and his multitude; and I will deliver him into thine hand.

ducam ad te in loco torrentis Cison Sisaram principem exercitus Iabin et currus eius atque omnem multitudinem et

tradam eos in manu tua 4:8 dixitque ad eam Barac si ue- 4:8 And Barak said unto her, If thou wilt go with me, then I will go: but if thou wilt not go with me, then I will not go.

nis mecum uadam si nolueris uenire non pergam 4:9 quae

4:9 And she said, I will surely go with thee: notwithstanding the journey that thou takest shall not be for thine honour; for the LORD shall sell Sisera into the hand of a woman. And Deborah arose, and went with Barak to Kedesh.

dixit ad eum ibo quidem tecum sed in hac uice tibi uictoria non reputabitur quia in manu mulieris tradetur Sisara surrexit itaque Debbora et perrexit cum Barac in Cedes 4:10 4:10 And Barak called Zebulun and Naphtali to Kedesh;

and he went up with ten thousand men at his feet: and Deborah went up with him.

qui accitis Zabulon et Nepthalim ascendit cum decem milibus pugnatorum habens Debboram in comitatu suo 4:11 4:11 Now Heber the Kenite, which was of the children

of Hobab the father in law of Moses, had severed himself from the Kenites, and pitched his tent unto the plain of

Aber autem Cineus recesserat quondam a ceteris Cineis fra- Zaanaim, which is by Kedesh. tribus suis filiis Obab cognati Mosi et tetenderat tabernacula usque ad uallem quae uocatur Sennim et erat iuxta Cedes 4:12 nuntiatumque est Sisarae quod ascendisset Barac filius 4:12 And they shewed Sisera that Barak the son of Abi-

noam was gone up to mount Tabor.

Abinoem in montem Thabor 4:13 et congregauit nongen- 4:13 And Sisera gathered together all his chariots, even nine hundred chariots of iron, and all the people that were with him, from Harosheth of the Gentiles unto the river of

tos falcatos currus omnemque exercitum de Aroseth gentium Kishon. ad torrentem Cison 4:14 dixitque Debbora ad Barac surge 4:14 And Deborah said unto Barak, Up; for this is the day

in which the LORD hath delivered Sisera into thine hand: is not the LORD gone out before thee? So Barak went

haec est enim dies in qua tradidit Dominus Sisaram in manus down from mount Tabor, and ten thousand men after him. tuas en ipse ductor est tuus descendit itaque Barac de monte Thabor et decem milia pugnatorum cum eo 4:15 perterruit- 4:15 And the LORD discomfited Sisera, and all his char-

iots, and all his host, with the edge of the sword before Barak; so that Sisera lighted down off his chariot, and fled

que Dominus Sisaram et omnes currus eius uniuersamque away on his feet. multitudinem in ore gladii ad conspectum Barac in tantum ut

Biblia Sacra Vulgata

4:16 But Barak pursued after the chariots, and after the Sisara de curru desiliens pedibus fugeret 4:16 et Barac per-

host, unto Harosheth of the Gentiles: and all the host of Sisera fell upon the edge of the sword; and there was not

a man left. sequeretur fugientes currus et exercitum usque ad Aroseth

gentium et omnis hostium multitudo usque ad internicionem

4:17 Howbeit Sisera fled away on his feet to the tent of Jael caderet 4:17 Sisara autem fugiens peruenit ad tentorium Ia-

the wife of Heber the Kenite: for there was peace between Jabin the king of Hazor and the house of Heber the Kenite.

hel uxoris Aber Cinei erat enim pax inter Iabin regem Asor

4:18 And Jael went out to meet Sisera, and said unto him, et domum Aber Cinei 4:18 egressa igitur Iahel in occursum

Turn in, my lord, turn in to me; fear not. And when he had turned in unto her into the tent, she covered him with

a mantle. Sisarae dixit ad eum intra ad me domine mi intra ne timeas 4:19 And he said unto her, Give me, I pray thee, a little qui ingressus tabernaculum eius et opertus ab ea pallio 4:19

water to drink; for I am thirsty. And she opened a bottle of milk, and gave him drink, and covered him.

dixit ad eam da mihi obsecro paululum aquae quia ualde sitio quae aperuit utrem lactis et dedit ei bibere et operuit il-

4:20 Again he said unto her, Stand in the door of the tent, lum 4:20 dixitque Sisara ad eam sta ante ostium tabernaand it shall be, when any man doth come and enquire of thee, and say, Is there any man here? that thou shalt say,

No. culi et cum uenerit aliquis interrogans te et dicens numquid 4:21 Then Jael Heber’s wife took a nail of the tent, and hic est aliquis respondebis nullus est 4:21 tulit itaque Iahel

took an hammer in her hand, and went softly unto him, and smote the nail into his temples, and fastened it into

the ground: for he was fast asleep and weary. So he died. uxor Aber clauum tabernaculi adsumens pariter malleum et ingressa abscondite et cum silentio posuit supra tempus capitis eius clauum percussumque malleo defixit in cerebrum usque ad terram qui soporem morti socians defecit et mor-

4:22 And, behold, as Barak pursued Sisera, Jael came out tuus est 4:22 et ecce Barac sequens Sisaram ueniebat egres-

to meet him, and said unto him, Come, and I will shew thee the man whom thou seekest. And when he came into her tent, behold, Sisera lay dead, and the nail was in his temples.

saque Iahel in occursum eius dixit ei ueni et ostendam tibi uirum quem quaeris qui cum intrasset ad eam uidit Sisaram

4:23 So God subdued on that day Jabin the king of Canaan iacentem mortuum et clauum infixum in tempore eius 4:23

before the children of Israel.

humiliauit ergo Deus in die illo Iabin regem Chanaan co-

4:24 And the hand of the children of Israel prospered, and ram filiis Israhel 4:24 qui crescebant cotidie et forti manu

prevailed against Jabin the king of Canaan, until they had destroyed Jabin king of Canaan.

opprimebant Iabin regem Chanaan donec delerent eum

5:1 Then sang Deborah and Barak the son of Abinoam on 5:1 cecineruntque Debbora et Barac filius Abinoem in die

that day, saying,

5:2 Praise ye the LORD for the avenging of Israel, when illo dicentes 5:2 qui sponte obtulistis de Israhel animas ue-

the people willingly offered themselves.

5:3 Hear, O ye kings; give ear, O ye princes; I, even I, will stras ad periculum benedicite Domino 5:3 audite reges per-

sing unto the LORD; I will sing praise to the LORD God of Israel.

cipite auribus principes ego sum ego sum quae Domino ca-

5:4 LORD, when thou wentest out of Seir, when thou nam psallam Domino Deo Israhel 5:4 Domine cum exires

marchedst out of the field of Edom, the earth trembled, and the heavens dropped, the clouds also dropped water.

de Seir et transires per regiones Edom terra mota est caeli-

5:5 The mountains melted from before the LORD, even que ac nubes stillauerunt aquis 5:5 montes fluxerunt a fa-

that Sinai from before the LORD God of Israel.

5:6 In the days of Shamgar the son of Anath, in the days cie Domini et Sinai a facie Domini Dei Israhel 5:6 in die-

of Jael, the highways were unoccupied, and the travellers walked through byways.

bus Samgar filii Anath in diebus Iahel quieuerunt semitae

Liber Iudicum

et qui ingrediebantur per eas ambulauerunt per calles deuios 5:7 cessauerunt fortes in Israhel et quieuerunt donec 5:7 The inhabitants of the villages ceased, they ceased in

Israel, until that I Deborah arose, that I arose a mother in Israel.

surgeret Debbora surgeret mater in Israhel 5:8 noua bella

5:8 They chose new gods; then was war in the gates: was there a shield or spear seen among forty thousand in Israel?

elegit Dominus et portas hostium ipse subuertit clypeus et hasta si apparuerint in quadraginta milibus Israhel 5:9 cor 5:9 My heart is toward the governors of Israel, that of-

fered themselves willingly among the people. Bless ye the LORD.

meum diligit principes Israhel qui propria uoluntate obtulistis uos discrimini benedicite Domino 5:10 qui ascenditis 5:10 Speak, ye that ride on white asses, ye that sit in judg-

ment, and walk by the way.

super nitentes asinos et sedetis in iudicio et ambulatis in uia loquimini 5:11 ubi conlisi sunt currus et hostium est suf- 5:11 They that are delivered from the noise of archers in

the places of drawing water, there shall they rehearse the righteous acts of the LORD, even the righteous acts toward the inhabitants of his villages in Israel: then shall the people of the LORD go down to the gates.

focatus exercitus ibi narrentur iustitiae Domini et clementia in fortes Israhel tunc descendit populus Domini ad portas et obtinuit principatum 5:12 surge surge Debbora surge surge 5:12 Awake, awake, Deborah: awake, awake, utter a song:

arise, Barak, and lead thy captivity captive, thou son of Abinoam.

et loquere canticum surge Barac et adprehende captiuos tuos fili Abinoem 5:13 saluatae sunt reliquiae populi Dominus 5:13 Then he made him that remaineth have dominion

over the nobles among the people: the LORD made me have dominion over the mighty.

in fortibus dimicauit 5:14 ex Ephraim deleuit eos in Ama-

5:14 Out of Ephraim was there a root of them against Amalek; after thee, Benjamin, among thy people; out of Machir came down governors, and out of Zebulun they that handle the pen of the writer.

lech et post eum ex Beniamin in populos tuos o Amalech de Machir principes descenderunt et de Zabulon qui exerci-

tum ducerent ad bellandum 5:15 duces Isachar fuere cum 5:15 And the princes of Issachar were with Deborah; even Issachar, and also Barak: he was sent on foot into the valley. For the divisions of Reuben there were great thoughts

Debbora et Barac uestigia sunt secuti qui quasi in praeceps of heart. ac baratrum se discrimini dedit diuiso contra se Ruben magnanimorum repperta contentio est 5:16 quare habitas in- 5:16 Why abodest thou among the sheepfolds, to hear the

bleatings of the flocks? For the divisions of Reuben there were great searchings of heart.

ter duos terminos ut audias sibilos gregum diuiso contra se Ruben magnanimorum repperta contentio est 5:17 Galaad 5:17 Gilead abode beyond Jordan: and why did Dan re-

main in ships? Asher continued on the sea shore, and abode in his breaches.

trans Iordanem quiescebat et Dan uacabat nauibus Aser habitabat in litore maris et in portibus morabatur 5:18 Zabu- 5:18 Zebulun and Naphtali were a people that jeoparded

their lives unto the death in the high places of the field.

lon uero et Nepthalim obtulerunt animas suas morti in regione Merome 5:19 uenerunt reges et pugnauerunt pugna- 5:19 The kings came and fought, then fought the kings of

Canaan in Taanach by the waters of Megiddo; they took no gain of money.

uerunt reges Chanaan in Thanach iuxta aquas Mageddo et tamen nihil tulere praedantes 5:20 de caelo dimicatum est 5:20 They fought from heaven; the stars in their courses

fought against Sisera.

contra eos stellae manentes in ordine et cursu suo aduersum Sisaram pugnauerunt 5:21 torrens Cison traxit cadauera eo- 5:21 The river of Kishon swept them away, that ancient

river, the river Kishon. O my soul, thou hast trodden down strength.

rum torrens Cadumim torrens Cison conculca anima mea robustos 5:22 ungulae equorum ceciderunt fugientibus im- 5:22 Then were the horsehoofs broken by the means of the

pransings, the pransings of their mighty ones.

Biblia Sacra Vulgata

5:23 Curse ye Meroz, said the angel of the LORD, curse petu et per praeceps ruentibus fortissimis hostium 5:23 ma-

ye bitterly the inhabitants thereof; because they came not to the help of the LORD, to the help of the LORD against

the mighty. ledicite terrae Meroz dixit angelus Domini maledicite habi-

tatoribus eius quia non uenerunt ad auxilium Domini in ad-

5:24 Blessed above women shall Jael the wife of Heber the iutorium fortissimorum eius 5:24 benedicta inter mulieres

Kenite be, blessed shall she be above women in the tent.

5:25 He asked water, and she gave him milk; she brought Iahel uxor Aber Cinei benedicatur in tabernaculo suo 5:25

forth butter in a lordly dish.

aquam petenti lac dedit et in fiala principum obtulit buty-

5:26 She put her hand to the nail, and her right hand to rum 5:26 sinistram manum misit ad clauum et dexteram ad

the workmen’s hammer; and with the hammer she smote Sisera, she smote off his head, when she had pierced and

stricken through his temples. fabrorum malleos percussitque Sisaram quaerens in capite 5:27 At her feet he bowed, he fell, he lay down: at her feet uulneri locum et tempus ualide perforans 5:27 inter pedes

he bowed, he fell: where he bowed, there he fell down dead.

eius ruit defecit et mortuus est ante pedes illius uoluebatur

5:28 The mother of Sisera looked out at a window, and et iacebat exanimis et miserabilis 5:28 per fenestram pro-

cried through the lattice, Why is his chariot so long in coming? why tarry the wheels of his chariots?

spiciens ululabat mater eius et de cenaculo loquebatur cur moratur regredi currus eius quare tardauerunt pedes quadri-

5:29 Her wise ladies answered her, yea, she returned an- garum illius 5:29 una sapientior ceteris uxoribus eius haec

swer to herself,

5:30 Have they not sped? have they not divided the prey; socrui uerba respondit 5:30 forsitan nunc diuidit spolia et

to every man a damsel or two; to Sisera a prey of divers colours, a prey of divers colours of needlework, of divers colours of needlework on both sides, meet for the necks of them that take the spoil?

pulcherrima feminarum eligitur ei uestes diuersorum colorum Sisarae traduntur in praedam et supellex uaria ad or-

5:31 So let all thine enemies perish, O LORD: but let them nanda colla congeritur 5:31 sic pereant omnes inimici tui

that love him be as the sun when he goeth forth in his might.

Domine qui autem diligunt te sicut sol in ortu suo splendet

5:32 And the land had rest forty years. ita rutilent 5:32 quieuitque terra per quadraginta annos

6:1 And the children of Israel did evil in the sight of the 6:1 fecerunt autem filii Israhel malum in conspectu Do-LORD: and the LORD delivered them into the hand of Midian seven years.

mini qui tradidit eos in manu Madian septem annis 6:2 et

6:2 And the hand of Midian prevailed against Israel: and because of the Midianites the children of Israel made them the dens which are in the mountains, and caves, and strong holds.

oppressi sunt ualde ab eis feceruntque sibi antra et speluncas

6:3 And so it was, when Israel had sown, that the Midian- in montibus et munitissima ad repugnandum loca 6:3 cum-

ites came up, and the Amalekites, and the children of the east, even they came up against them;

que seuisset Israhel ascendebat Madian et Amalech et ceteri

6:4 And they encamped against them, and destroyed the orientalium nationum 6:4 et apud eos figentes tentoria sic-

increase of the earth, till thou come unto Gaza, and left no sustenance for Israel, neither sheep, nor ox, nor ass.

ut erant in herbis cuncta uastabant usque ad introitum Gazae nihilque omnino ad uitam pertinens relinquebant in Israhel

6:5 For they came up with their cattle and their tents, and non oues non boues non asinos 6:5 ipsi enim et uniuersi they came as grasshoppers for multitude; for both they and their camels were without number: and they entered into

the land to destroy it. greges eorum ueniebant cum tabernaculis et instar lucusta-

rum uniuersa conplebant innumera multitudo hominum et

6:6 And Israel was greatly impoverished because of the camelorum quicquid tetigerant deuastantes 6:6 humiliatus-

Midianites; and the children of Israel cried unto the LORD.

que est Israhel ualde in conspectu Madian 6:7 et clamauit

6:7 And it came to pass, when the children of Israel cried unto the LORD because of the Midianites,

Liber Iudicum

ad Dominum postulans auxilium contra Madianitas 6:8 qui 6:8 That the LORD sent a prophet unto the children of

Israel, which said unto them, Thus saith the LORD God of Israel, I brought you up from Egypt, and brought you

misit ad eos uirum prophetam et locutus est haec dicit Do- forth out of the house of bondage; minus Deus Israhel ego uos feci conscendere de Aegypto et eduxi de domo seruitutis 6:9 et liberaui de manu Aegyptio- 6:9 And I delivered you out of the hand of the Egyptians,

and out of the hand of all that oppressed you, and drave them out from before you, and gave you their land;

rum et omnium inimicorum qui adfligebant uos eiecique eos ad introitum uestrum et tradidi uobis terram eorum 6:10 et 6:10 And I said unto you, I am the LORD your God; fear

not the gods of the Amorites, in whose land ye dwell: but ye have not obeyed my voice.

dixi ego Dominus Deus uester ne timeatis deos Amorreorum in quorum terra habitatis et noluistis audire uocem meam 6:11 uenit autem angelus Domini et sedit sub quercu quae 6:11 And there came an angel of the LORD, and sat under

an oak which was in Ophrah, that pertained unto Joash the Abiezrite: and his son Gideon threshed wheat by the

erat in Ephra et pertinebat ad Ioas patrem familiae Ezri cum- winepress, to hide it from the Midianites. que Gedeon filius eius excuteret atque purgaret frumenta in torculari ut fugeret Madian 6:12 apparuit ei et ait Dominus 6:12 And the angel of the LORD appeared unto him, and

said unto him, The LORD is with thee, thou mighty man of valour.

tecum uirorum fortissime 6:13 dixitque ei Gedeon obsecro

6:13 And Gideon said unto him, Oh my Lord, if the LORD be with us, why then is all this befallen us? and where be all his miracles which our fathers told us of, saying, Did not the LORD bring us up from Egypt? but now the LORD hath forsaken us, and delivered us into the hands of the Midianites.

Domine si Dominus nobiscum est cur adprehenderunt nos haec omnia ubi sunt mirabilia eius quae narrauerunt patres nostri atque dixerunt de Aegypto eduxit nos Dominus nunc autem dereliquit nos et tradidit in manibus Madian 6:14 re- 6:14 And the LORD looked upon him, and said, Go in this

thy might, and thou shalt save Israel from the hand of the Midianites: have not I sent thee?

spexitque ad eum Dominus et ait uade in hac fortitudine tua et liberabis Israhel de manu Madian scito quod miserim te 6:15 qui respondens ait obsecro Domine mi in quo liberabo 6:15 And he said unto him, Oh my Lord, wherewith shall

I save Israel? behold, my family is poor in Manasseh, and I am the least in my father’s house.

Israhel ecce familia mea infima est in Manasse et ego minimus in domo patris mei 6:16 dixitque ei Dominus ego ero 6:16 And the LORD said unto him, Surely I will be with

thee, and thou shalt smite the Midianites as one man.

tecum et percuties Madian quasi unum uirum 6:17 et ille si 6:17 And he said unto him, If now I have found grace in

thy sight, then shew me a sign that thou talkest with me.

inueni inquit gratiam coram te da mihi signum quod tu sis

qui loquaris ad me 6:18 ne recedas hinc donec reuertar ad 6:18 Depart not hence, I pray thee, until I come unto thee, and bring forth my present, and set it before thee. And he said, I will tarry until thou come again.

te portans sacrificium et offerens tibi qui respondit ego praestolabor aduentum tuum 6:19 ingressus est itaque Gedeon 6:19 And Gideon went in, and made ready a kid, and un-

leavened cakes of an ephah of flour: the flesh he put in a basket, and he put the broth in a pot, and brought it out

et coxit hedum et de farinae modio azymos panes carnes- unto him under the oak, and presented it. que ponens in canistro et ius carnium mittens in ollam tulit omnia sub quercum et obtulit ei 6:20 cui dixit angelus Do- 6:20 And the angel of God said unto him, Take the flesh

and the unleavened cakes, and lay them upon this rock, and pour out the broth. And he did so.

mini tolle carnes et panes azymos et pone super petram illam et ius desuper funde cumque fecisset ita 6:21 extendit an- 6:21 Then the angel of the LORD put forth the end of the

staff that was in his hand, and touched the flesh and the unleavened cakes; and there rose up fire out of the rock, and consumed the flesh and the unleavened cakes. Then the angel of the LORD departed out of his sight.

gelus Domini summitatem uirgae quam tenebat in manu et

Biblia Sacra Vulgata

tetigit carnes et azymos panes ascenditque ignis de petra et carnes azymosque consumpsit angelus autem Domini eua-

6:22 And when Gideon perceived that he was an angel of nuit ex oculis eius 6:22 uidensque Gedeon quod esset an-

the LORD, Gideon said, Alas, O LORD God! for because I have seen an angel of the LORD face to face.

gelus Domini ait heu mihi Domine Deus quia uidi angelum

6:23 And the LORD said unto him, Peace be unto thee; Domini facie ad faciem 6:23 dixitque ei Dominus pax te-

fear not: thou shalt not die. 6:24 Then Gideon built an altar there unto the LORD, and cum ne timeas non morieris 6:24 aedificauit ergo ibi Gecalled it Jehovahshalom: unto this day it is yet in Ophrah of the Abiezrites.

deon altare Domino uocauitque illud Domini pax usque in praesentem diem cum adhuc esset in Ephra quae est fami-

6:25 And it came to pass the same night, that the LORD liae Ezri 6:25 nocte illa dixit Dominus ad eum tolle tau-

said unto him, Take thy father’s young bullock, even the second bullock of seven years old, and throw down the altar of Baal that thy father hath, and cut down the grove that is by it:

rum patris tui et alterum taurum annorum septem destruesque aram Baal quae est patris tui et nemus quod circa aram

6:26 And build an altar unto the LORD thy God upon the est succide 6:26 et aedificabis altare Domino Deo tuo in

top of this rock, in the ordered place, and take the second bullock, and offer a burnt sacrifice with the wood of the

grove which thou shalt cut down. summitate petrae huius super quam sacrificium ante posuisti

tollesque taurum secundum et offeres holocaustum super li-

6:27 Then Gideon took ten men of his servants, and did as gnorum struem quae de nemore succideris 6:27 adsumptis

the LORD had said unto him: and so it was, because he feared his father’s household, and the men of the city, that

he could not do it by day, that he did it by night. igitur Gedeon decem uiris de seruis suis fecit sicut praeceperat Dominus timens autem domum patris sui et homines illius ciuitatis per diem facere noluit sed omnia nocte conple-

6:28 And when the men of the city arose early in the morn- uit 6:28 cumque surrexissent uiri oppidi eius mane uiderunt

ing, behold, the altar of Baal was cast down, and the grove was cut down that was by it, and the second bullock was

offered upon the altar that was built. destructam aram Baal lucumque succisum et taurum alte-6:29 And they said one to another, Who hath done this rum inpositum super altare quod tunc aedificatum erat 6:29

thing? And when they enquired and asked, they said, Gideon the son of Joash hath done this thing.

dixeruntque ad inuicem quis hoc fecit cumque perquirerent auctorem facti dictum est Gedeon filius Ioas fecit haec om-

6:30 Then the men of the city said unto Joash, Bring out nia 6:30 et dixerunt ad Ioas produc filium tuum ut moriatur

thy son, that he may die: because he hath cast down the altar of Baal, and because he hath cut down the grove that

was by it. quia destruxit aram Baal et succidit nemus 6:31 quibus ille

6:31 And Joash said unto all that stood against him, Will ye plead for Baal? will ye save him? he that will plead for him, let him be put to death whilst it is yet morning: if he be a god, let him plead for himself, because one hath cast down his altar.

respondit numquid ultores estis Baal et pugnatis pro eo qui aduersarius eius est moriatur antequam lux crastina ueniat

6:32 Therefore on that day he called him Jerubbaal, say- si deus est uindicet se de eo qui suffodit aram eius 6:32 ing, Let Baal plead against him, because he hath thrown down his altar.

ex illo die uocatus est Gedeon Hierobbaal eo quod dixisset

6:33 Then all the Midianites and the Amalekites and the Ioas ulciscatur se de eo Baal qui suffodit altare eius 6:33

children of the east were gathered together, and went over, and pitched in the valley of Jezreel.

igitur omnis Madian et Amalech et orientales populi congregati sunt simul et transeuntes Iordanem castrametati sunt

6:34 But the Spirit of the LORD came upon Gideon, and in ualle Iezrahel 6:34 spiritus autem Domini induit Gedeon

he blew a trumpet; and Abiezer was gathered after him.

Liber Iudicum

qui clangens bucina conuocauit domum Abiezer ut sequeretur 6:35 misitque nuntios in uniuersum Manassen qui et 6:35 And he sent messengers throughout all Manasseh;

who also was gathered after him: and he sent messengers unto Asher, and unto Zebulun, and unto Naphtali; and they

ipse secutus est eum et alios nuntios in Aser et Zabulon et came up to meet them. Nepthalim qui occurrerunt ei 6:36 dixitque Gedeon ad Do- 6:36 And Gideon said unto God, If thou wilt save Israel

by mine hand, as thou hast said,

minum si saluum facis per manum meam Israhel sicut locutus es 6:37 ponam uellus hoc lanae in area si ros in solo 6:37 Behold, I will put a fleece of wool in the floor; and

if the dew be on the fleece only, and it be dry upon all the earth beside, then shall I know that thou wilt save Israel by

uellere fuerit et in omni terra siccitas sciam quod per manum mine hand, as thou hast said. meam sicut locutus es liberabis Israhel 6:38 factumque est 6:38 And it was so: for he rose up early on the morrow,

and thrust the fleece together, and wringed the dew out of the fleece, a bowl full of water.

ita et de nocte consurgens expresso uellere concam rore conpleuit 6:39 dixitque rursus ad Dominum ne irascatur furor 6:39 And Gideon said unto God, Let not thine anger be hot

against me, and I will speak but this once: let me prove, I pray thee, but this once with the fleece; let it now be dry only upon the fleece, and upon all the ground let there be dew.

tuus contra me si adhuc semel temptauero signum quaerens in uellere oro ut solum uellus siccum sit et omnis terra rore madens 6:40 fecitque Dominus nocte illa ut postulauerat et 6:40 And God did so that night: for it was dry upon the

fleece only, and there was dew on all the ground.

fuit siccitas in solo uellere et ros in omni terra

7:1 igitur Hierobbaal qui est et Gedeon de nocte consur- 7:1 Then Jerubbaal, who is Gideon, and all the people that were with him, rose up early, and pitched beside the well of Harod: so that the host of the Midianites were on the

gens et omnis populus cum eo uenit ad fontem qui uoca- north side of them, by the hill of Moreh, in the valley.

tur Arad erant autem castra Madian in ualle ad septentrionalem plagam collis Excelsi 7:2 dixitque Dominus ad Ge- 7:2 And the LORD said unto Gideon, The people that are

with thee are too many for me to give the Midianites into their hands, lest Israel vaunt themselves against me, say-

deon multus tecum est populus nec tradetur Madian in ma- ing, Mine own hand hath saved me. nus eius ne glorietur contra me Israhel et dicat meis uiribus liberatus sum 7:3 loquere ad populum et cunctis audienti- 7:3 Now therefore go to, proclaim in the ears of the people,

saying, Whosoever is fearful and afraid, let him return and depart early from mount Gilead. And there returned of the people twenty and two thousand; and there remained ten thousand.

bus praedica qui formidolosus et timidus est reuertatur recesseruntque de monte Galaad et reuersa sunt ex populo uiginti duo milia uirorum et tantum decem milia remanserunt 7:4 dixitque Dominus ad Gedeon adhuc populus multus est 7:4 And the LORD said unto Gideon, The people are yet

too many; bring them down unto the water, and I will try them for thee there: and it shall be, that of whom I say unto thee, This shall go with thee, the same shall go with thee; and of whomsoever I say unto thee, This shall not go with thee, the same shall not go.

duc eos ad aquas et ibi probabo illos et de quo dixero tibi ut tecum uadat ipse pergat quem ire prohibuero reuertatur 7:5 cumque descendisset populus ad aquas dixit Dominus ad 7:5 So he brought down the people unto the water: and

the LORD said unto Gideon, Every one that lappeth of the water with his tongue, as a dog lappeth, him shalt thou set by himself; likewise every one that boweth down upon his knees to drink.

Gedeon qui lingua lambuerint aquas sicut solent canes lambere separabis eos seorsum qui autem curuatis genibus biberint in altera parte erunt 7:6 fuit itaque numerus eorum qui 7:6 And the number of them that lapped, putting their hand

to their mouth, were three hundred men: but all the rest of the people bowed down upon their knees to drink water.

manu ad os proiciente aquas lambuerant trecenti uiri omnis autem reliqua multitudo flexo poplite biberat 7:7 et ait 7:7 And the LORD said unto Gideon, By the three hundred

men that lapped will I save you, and deliver the Midianites into thine hand: and let all the other people go every man unto his place.

Biblia Sacra Vulgata

Dominus ad Gedeon in trecentis uiris qui lambuerunt aquas liberabo uos et tradam Madian in manu tua omnis autem reli-

7:8 So the people took victuals in their hand, and their qua multitudo reuertatur in locum suum 7:8 sumptis itaque

trumpets: and he sent all the rest of Israel every man unto his tent, and retained those three hundred men: and the

host of Midian was beneath him in the valley. pro numero cibariis et tubis omnem reliquam multitudinem abire praecepit ad tabernacula sua et ipse cum trecentis uiris se certamini dedit castra autem Madian erant subter in

7:9 And it came to pass the same night, that the LORD ualle 7:9 eadem nocte dixit Dominus ad eum surge et de-

said unto him, Arise, get thee down unto the host; for I have delivered it into thine hand.

scende in castra quia tradidi eos in manu tua 7:10 sin autem

7:10 But if thou fear to go down, go thou with Phurah thy servant down to the host:

7:11 And thou shalt hear what they say; and afterward solus ire formidas descendat tecum Phara puer tuus 7:11 et

shall thine hands be strengthened to go down unto the host. Then went he down with Phurah his servant unto the out-

side of the armed men that were in the host. cum audieris quid loquantur tunc confortabuntur manus tuae et securior ad hostium castra descendes descendit ergo ipse et Phara puer eius in partem castrorum ubi erant armatorum

7:12 And the Midianites and the Amalekites and all the uigiliae 7:12 Madian autem et Amalech et omnes orienta-

children of the east lay along in the valley like grasshoppers for multitude; and their camels were without number,

as the sand by the sea side for multitude. les populi fusi iacebant in ualle ut lucustarum multitudo ca-

meli quoque innumerabiles erant sicut harena quae iacet in

7:13 And when Gideon was come, behold, there was a litoribus maris 7:13 cumque uenisset Gedeon narrabat ali-

man that told a dream unto his fellow, and said, Behold, I dreamed a dream, and, lo, a cake of barley bread tumbled into the host of Midian, and came unto a tent, and smote it that it fell, and overturned it, that the tent lay along.

quis somnium proximo suo et in hunc modum referebat quod uiderat uidi somnium et uidebatur mihi quasi subcinericius panis ex hordeo uolui et in Madian castra descendere cumque peruenisset ad tabernaculum percussit illud atque sub-

7:14 And his fellow answered and said, This is nothing uertit et terrae funditus coaequauit 7:14 respondit is cui lo-

else save the sword of Gideon the son of Joash, a man of Israel: for into his hand hath God delivered Midian, and

all the host. quebatur non est hoc aliud nisi gladius Gedeonis filii Ioas

uiri Israhelitae tradidit Deus in manu eius Madian et om-

7:15 And it was so, when Gideon heard the telling of the nia castra eius 7:15 cumque audisset Gedeon somnium et

dream, and the interpretation thereof, that he worshipped, and returned into the host of Israel, and said, Arise; for the

LORD hath delivered into your hand the host of Midian. interpretationem eius adorauit et reuersus ad castra Israhel

ait surgite tradidit enim Dominus in manus nostras castra

7:16 And he divided the three hundred men into three Madian 7:16 diuisitque trecentos uiros in tres partes et de-

companies, and he put a trumpet in every man’s hand, with empty pitchers, and lamps within the pitchers.

dit tubas in manibus eorum lagoenasque uacuas ac lampadas

7:17 And he said unto them, Look on me, and do likewise: in medio lagoenarum 7:17 et dixit ad eos quod me facere and, behold, when I come to the outside of the camp, it shall be that, as I do, so shall ye do.

uideritis hoc facite ingrediar partem castrorum et quod fe-

7:18 When I blow with a trumpet, I and all that are with cero sectamini 7:18 quando personauerit tuba in manu mea

me, then blow ye the trumpets also on every side of all the camp, and say, The sword of the LORD, and of Gideon.

uos quoque per castrorum circuitum clangite et conclamate

7:19 So Gideon, and the hundred men that were with him, Domino et Gedeoni 7:19 ingressusque est Gedeon et tre-

came unto the outside of the camp in the beginning of the middle watch; and they had but newly set the watch: and they blew the trumpets, and brake the pitchers that were in their hands.

Liber Iudicum

centi uiri qui erant cum eo in parte castrorum incipientibus uigiliis noctis mediae et custodibus suscitatis coeperunt bucinis clangere et conplodere inter se lagoenas 7:20 cumque 7:20 And the three companies blew the trumpets, and

brake the pitchers, and held the lamps in their left hands, and the trumpets in their right hands to blow withal: and

per gyrum castrorum in tribus personarent locis et hydrias they cried, The sword of the LORD, and of Gideon. confregissent tenuerunt sinistris manibus lampadas et dextris sonantes tubas clamaueruntque gladius Domini et Gedeonis 7:21 stantes singuli in loco suo per circuitum castrorum ho- 7:21 And they stood every man in his place round about

the camp; and all the host ran, and cried, and fled.

stilium omnia itaque castra turbata sunt et uociferantes ululantesque fugerunt 7:22 et nihilominus insistebant trecenti 7:22 And the three hundred blew the trumpets, and the

LORD set every man’s sword against his fellow, even throughout all the host: and the host fled to Bethshittah in Zererath, and to the border of Abelmeholah, unto Tabbath.

uiri bucinis personantes inmisitque Dominus gladium in omnibus castris et mutua se caede truncabant 7:23 fugientes 7:23 And the men of Israel gathered themselves together

out of Naphtali, and out of Asher, and out of all Manasseh, and pursued after the Midianites.

usque Bethseta et crepidinem Abelmeula in Tebbath conclamantes autem uiri Israhel de Nepthali et Aser et omni Manasse persequebantur Madian 7:24 misitque Gedeon nun- 7:24 And Gideon sent messengers throughout all mount

Ephraim, saying, come down against the Midianites, and take before them the waters unto Bethbarah and Jordan. Then all the men of Ephraim gathered themselves together, and took the waters unto Bethbarah and Jordan.

tios in omnem montem Ephraim dicens descendite in occursum Madian et occupate aquas usque Bethbera atque Iordanem clamauitque omnis Ephraim et praeoccupauit aquas atque Iordanem usque Bethbera 7:25 adprehensosque duos 7:25 And they took two princes of the Midianites, Oreb

and Zeeb; and they slew Oreb upon the rock Oreb, and Zeeb they slew at the winepress of Zeeb, and pursued Midian, and brought the heads of Oreb and Zeeb to Gideon on the other side Jordan.

uiros Madian Oreb et Zeb interfecit Oreb in petra Oreb Zeb uero in torculari Zeb et persecuti sunt Madian capita Oreb et Zeb portantes ad Gedeon trans fluenta Iordanis

8:1 dixeruntque ad eum uiri Ephraim quid est hoc quod 8:1 And the men of Ephraim said unto him, Why hast thou

served us thus, that thou calledst us not, when thou wentest to fight with the Midianites? And they did chide with him

facere uoluisti ut non nos uocares cum ad pugnam perge- sharply. res contra Madian iurgantes fortiter et prope uim inferentes 8:2 quibus ille respondit quid enim tale facere potui quale 8:2 And he said unto them, What have I done now in

comparison of you? Is not the gleaning of the grapes of Ephraim better than the vintage of Abiezer?

uos fecistis nonne melior est racemus Ephraim uindemiis Abiezer 8:3 in manus uestras tradidit Dominus principes 8:3 God hath delivered into your hands the princes of Mid-

ian, Oreb and Zeeb: and what was I able to do in comparison of you? Then their anger was abated toward him,

Madian Oreb et Zeb quid tale facere potui quale uos fe- when he had said that. cistis quod cum locutus esset requieuit spiritus eorum quo tumebant contra eum 8:4 cumque uenisset Gedeon ad Ior- 8:4 And Gideon came to Jordan, and passed over, he, and

the three hundred men that were with him, faint, yet pursuing them.

danem transiuit eum cum trecentis uiris qui secum erant et prae lassitudine fugientes persequi non poterant 8:5 dixit- 8:5 And he said unto the men of Succoth, Give, I pray you,

loaves of bread unto the people that follow me; for they be faint, and I am pursuing after Zebah and Zalmunna, kings

que ad uiros Soccoth date obsecro panes populo qui me- of Midian.

Biblia Sacra Vulgata

cum est quia ualde defecerunt ut possimus persequi Zebee

8:6 And the princes of Succoth said, Are the hands of Ze- et Salmana reges Madian 8:6 responderunt principes Soc-

bah and Zalmunna now in thine hand, that we should give bread unto thine army?

coth forsitan palmae manuum Zebee et Salmana in manu tua

8:7 And Gideon said, Therefore when the LORD hath de- sunt et idcirco postulas ut demus exercitui tuo panes 8:7

livered Zebah and Zalmunna into mine hand, then I will tear your flesh with the thorns of the wilderness and with

briers. quibus ille ait cum ergo tradiderit Dominus Zebee et Sal-

mana in manus meas conteram carnes uestras cum spinis tri-

8:8 And he went up thence to Penuel, and spake unto them bulisque deserti 8:8 et inde conscendens uenit in Phanuhel

likewise: and the men of Penuel answered him as the men of Succoth had answered him.

locutusque est ad uiros eius loci similia cui et illi respon-

8:9 And he spake also unto the men of Penuel, saying, derunt sicut responderant uiri Soccoth 8:9 dixit itaque et

When I come again in peace, I will break down this tower.

eis cum reuersus fuero uictor in pace destruam turrem hanc

8:10 Now Zebah and Zalmunna were in Karkor, and their 8:10 Zebee autem et Salmana requiescebant cum omni ex-

hosts with them, about fifteen thousand men, all that were left of all the hosts of the children of the east: for there fell

an hundred and twenty thousand men that drew sword. ercitu suo quindecim milia enim uiri remanserant ex om-

nibus turmis orientalium populorum caesis centum uiginti

8:11 And Gideon went up by the way of them that dwelt milibus bellatorum et educentium gladium 8:11 ascendens-

in tents on the east of Nobah and Jogbehah, and smote the host; for the host was secure.

que Gedeon per uiam eorum qui in tabernaculis morabantur ad orientalem partem Nobee et Iecbaa percussit castra hos-

8:12 And when Zebah and Zalmunna fled, he pursued af- tium qui securi erant et nihil aduersi suspicabantur 8:12 fu-

ter them, and took the two kings of Midian, Zebah and Zalmunna, and discomfited all the host.

geruntque Zebee et Salmana quos persequens Gedeon con-

8:13 And Gideon the son of Joash returned from battle prehendit turbato omni exercitu eorum 8:13 reuertensque

before the sun was up,

8:14 And caught a young man of the men of Succoth, and de bello ante solis ortum 8:14 adprehendit puerum de uiris

enquired of him: and he described unto him the princes of Succoth, and the elders thereof, even threescore and sev-

enteen men. Soccoth interrogauitque eum nomina principum et seniorum 8:15 And he came unto the men of Succoth, and said, Soccoth et descripsit septuaginta septem uiros 8:15 uenit-

Behold Zebah and Zalmunna, with whom ye did upbraid me, saying, Are the hands of Zebah and Zalmunna now in thine hand, that we should give bread unto thy men that are weary?

que ad Soccoth et dixit eis en Zebee et Salmana super quibus exprobrastis mihi dicentes forsitan manus Zebee et Salmana in manibus tuis sunt et idcirco postulas ut demus uiris

8:16 And he took the elders of the city, and thorns of the qui lassi sunt et defecerunt panes 8:16 tulit ergo seniores

wilderness and briers, and with them he taught the men of Succoth.

ciuitatis et spinas deserti ac tribulos et contriuit cum eis at-

8:17 And he beat down the tower of Penuel, and slew the que comminuit uiros Soccoth 8:17 turrem quoque Phanuhel

men of the city.

8:18 Then said he unto Zebah and Zalmunna, What man- subuertit occisis habitatoribus ciuitatis 8:18 dixitque ad Ze-

ner of men were they whom ye slew at Tabor? And they answered, As thou art, so were they; each one resembled

the children of a king. bee et Salmana quales fuerunt uiri quos occidistis in Thabor

qui responderunt similes tui et unus ex eis quasi filius re-

8:19 And he said, They were my brethren, even the sons gis 8:19 quibus ille ait fratres mei fuerunt filii matris meae

of my mother: as the LORD liveth, if ye had saved them alive, I would not slay you.

uiuit Dominus si seruassetis eos non uos occiderem 8:20

8:20 And he said unto Jether his firstborn, Up, and slay them. But the youth drew not his sword: for he feared, because he was yet a youth.

Liber Iudicum

dixitque Ietther primogenito suo surge et interfice eos qui non eduxit gladium timebat enim quia adhuc puer erat 8:21 8:21 Then Zebah and Zalmunna said, Rise thou, and fall

upon us: for as the man is, so is his strength. And Gideon arose, and slew Zebah and Zalmunna, and took away the

dixeruntque Zebee et Salmana tu surge et inrue in nos quia ornaments that were on their camels’ necks. iuxta aetatem robur est hominis surrexit Gedeon et interfecit Zebee et Salmana et tulit ornamenta ac bullas quibus colla regalium camelorum decorari solent 8:22 dixeruntque om- 8:22 Then the men of Israel said unto Gideon, Rule thou

over us, both thou, and thy son, and thy son’s son also: for thou hast delivered us from the hand of Midian.

nes uiri Israhel ad Gedeon dominare nostri tu et filius tuus et filius filii tui quia liberasti nos de manu Madian 8:23 quibus 8:23 And Gideon said unto them, I will not rule over you,

neither shall my son rule over you: the LORD shall rule over you.

ille ait non dominabor uestri nec dominabitur in uos filius meus sed dominabitur Dominus 8:24 dixitque ad eos unam 8:24 And Gideon said unto them, I would desire a request

of you, that ye would give me every man the earrings of his prey. (For they had golden earrings, because they were

petitionem postulo a uobis date mihi inaures ex praeda ues- Ishmaelites.) tra inaures enim aureas Ismahelitae habere consuerant 8:25 8:25 And they answered, We will willingly give them.

And they spread a garment, and did cast therein every man the earrings of his prey.

qui responderunt libentissime dabimus expandentesque super terram pallium proiecerunt in eo inaures de praeda 8:26 8:26 And the weight of the golden earrings that he re-

quested was a thousand and seven hundred shekels of gold; beside ornaments, and collars, and purple raiment that was on the kings of Midian, and beside the chains that were about their camels’ necks.

et fuit pondus postulatarum inaurium mille septingenti auri sicli absque ornamentis et monilibus et ueste purpurea quibus Madian reges uti soliti erant et praeter torques aureos camelorum 8:27 fecitque ex eo Gedeon ephod et posuit il- 8:27 And Gideon made an ephod thereof, and put it in his

city, even in Ophrah: and all Israel went thither a whoring after it: which thing became a snare unto Gideon, and to

lud in ciuitate sua Ephra fornicatusque est omnis Israhel in his house. eo et factum est Gedeoni et omni domui eius in ruinam 8:28 8:28 Thus was Midian subdued before the children of Is-

rael, so that they lifted up their heads no more. And the country was in quietness forty years in the days of Gideon.

humiliatus est autem Madian coram filiis Israhel nec potuerunt ultra eleuare ceruices sed quieuit terra per quadraginta annos quibus praefuit Gedeon 8:29 abiit itaque Hierobbaal 8:29 And Jerubbaal the son of Joash went and dwelt in his

own house.

filius Ioas et habitauit in domo sua 8:30 habuitque septua- 8:30 And Gideon had threescore and ten sons of his body

begotten: for he had many wives.

ginta filios qui egressi sunt de femore eius eo quod plures haberet uxores 8:31 concubina autem illius quam habebat 8:31 And his concubine that was in Shechem, she also bare

him a son, whose name he called Abimelech.

in Sychem genuit ei filium nomine Abimelech 8:32 mor- 8:32 And Gideon the son of Joash died in a good old age,

and was buried in the sepulchre of Joash his father, in Ophrah of the Abiezrites.

tuusque est Gedeon filius Ioas in senectute bona et sepultus in sepulchro Ioas patris sui in Ephra de familia Ezri 8:33 8:33 And it came to pass, as soon as Gideon was dead, that

the children of Israel turned again, and went a whoring after Baalim, and made Baalberith their god.

postquam autem mortuus est Gedeon auersi sunt filii Israhel et fornicati cum Baalim percusseruntque cum Baal foedus

ut esset eis in deum 8:34 nec recordati sunt Domini Dei 8:34 And the children of Israel remembered not the LORD their God, who had delivered them out of the hands of all their enemies on every side:

sui qui eruit eos de manu omnium inimicorum suorum per

Biblia Sacra Vulgata

8:35 Neither shewed they kindness to the house of Jerub- circuitum 8:35 nec fecerunt misericordiam cum domo Hie-

baal, namely, Gideon, according to all the goodness which he had shewed unto Israel.

robbaal Gedeon iuxta omnia bona quae fecerat Israheli

9:1 And Abimelech the son of Jerubbaal went to Shechem 9:1 abiit autem Abimelech filius Hierobbaal in Sychem ad

unto his mother’s brethren, and communed with them, and with all the family of the house of his mother’s father, say-

ing, fratres matris suae et locutus est ad eos et ad omnem cog-9:2 Speak, I pray you, in the ears of all the men of nationem domus patris matris suae dicens 9:2 loquimini ad

Shechem, Whether is better for you, either that all the sons of Jerubbaal, which are threescore and ten persons, reign over you, or that one reign over you? remember also that I am your bone and your flesh.

omnes uiros Sychem quid uobis est melius ut dominentur uestri septuaginta uiri omnes filii Hierobbaal an ut dominetur uobis unus uir simulque considerate quia os uestrum et

9:3 And his mother’s brethren spake of him in the ears of caro uestra sum 9:3 locutique sunt fratres matris eius de

all the men of Shechem all these words: and their hearts inclined to follow Abimelech; for they said, He is our

brother. eo ad omnes uiros Sychem uniuersos sermones istos et in-

clinauerunt cor eorum post Abimelech dicentes frater no-

9:4 And they gave him threescore and ten pieces of silver ster est 9:4 dederuntque illi septuaginta pondo argenti de

out of the house of Baalberith, wherewith Abimelech hired vain and light persons, which followed him.

fano Baalbrith qui conduxit sibi ex eo uiros inopes et ua-

9:5 And he went unto his father’s house at Ophrah, and gos secutique sunt eum 9:5 et uenit in domum patris sui

slew his brethren the sons of Jerubbaal, being threescore and ten persons, upon one stone: notwithstanding yet Jotham the youngest son of Jerubbaal was left; for he hid himself.

Ephra et occidit fratres suos filios Hierobbaal septuaginta uiros super lapidem unum remansitque Ioatham filius Hie-

9:6 And all the men of Shechem gathered together, and all robbaal minimus et absconditus est 9:6 congregati sunt au-

the house of Millo, and went, and made Abimelech king, by the plain of the pillar that was in Shechem.

tem omnes uiri Sychem et uniuersae familiae urbis Mello abieruntque et constituerunt regem Abimelech iuxta quer-

9:7 And when they told it to Jotham, he went and stood cum quae stabat in Sychem 9:7 quod cum nuntiatum esset

in the top of mount Gerizim, and lifted up his voice, and cried, and said unto them, Hearken unto me, ye men of

Shechem, that God may hearken unto you. Ioatham iuit et stetit in uertice montis Garizim eleuataque

uoce clamauit et dixit audite me uiri Sychem ita audiat uos

9:8 The trees went forth on a time to anoint a king over Deus 9:8 ierunt ligna ut unguerent super se regem dixe-

them; and they said unto the olive tree, Reign thou over us.

runtque oliuae impera nobis 9:9 quae respondit numquid

9:9 But the olive tree said unto them, Should I leave my fatness, wherewith by me they honour God and man, and go to be promoted over the trees?

possum deserere pinguedinem meam qua et dii utuntur et

9:10 And the trees said to the fig tree, Come thou, and homines et uenire ut inter ligna promouear 9:10 dixerunt-

reign over us.

que ligna ad arborem ficum ueni et super nos regnum accipe

9:11 But the fig tree said unto them, Should I forsake my 9:11 quae respondit eis numquid possum deserere dulcedi-

sweetness, and my good fruit, and go to be promoted over the trees?

nem meam fructusque suauissimos et ire ut inter cetera ligna

9:12 Then said the trees unto the vine, Come thou, and commouear 9:12 locuta sunt quoque ligna ad uitem ueni et

reign over us. 9:13 And the vine said unto them, Should I leave my wine, impera nobis 9:13 quae respondit numquid possum desewhich cheereth God and man, and go to be promoted over the trees?

rere uinum meum quod laetificat Deum et homines et inter

9:14 Then said all the trees unto the bramble, Come thou, ligna cetera commoueri 9:14 dixeruntque omnia ligna ad and reign over us.

Liber Iudicum

ramnum ueni et impera super nos 9:15 quae respondit eis 9:15 And the bramble said unto the trees, If in truth ye

anoint me king over you, then come and put your trust in my shadow: and if not, let fire come out of the bramble,

si uere me regem uobis constituitis uenite et sub mea umbra and devour the cedars of Lebanon. requiescite sin autem non uultis egrediatur ignis de ramno et deuoret cedros Libani 9:16 nunc igitur si recte et abs- 9:16 Now therefore, if ye have done truly and sincerely,

in that ye have made Abimelech king, and if ye have dealt well with Jerubbaal and his house, and have done unto him

que peccato constituistis super uos regem Abimelech et bene according to the deserving of his hands; egistis cum Hierobbaal et cum domo eius et reddidistis ui-

cem beneficiis eius qui pugnauit pro uobis 9:17 et animam 9:17 (For my father fought for you, and adventured his life

far, and delivered you out of the hand of Midian:

suam dedit periculis ut erueret uos de manu Madian 9:18 9:18 And ye are risen up against my father’s house this

day, and have slain his sons, threescore and ten persons, upon one stone, and have made Abimelech, the son of his maidservant, king over the men of Shechem, because he is your brother;)

qui nunc surrexistis contra domum patris mei et interfecistis filios eius septuaginta uiros super unum lapidem et constituistis regem Abimelech filium ancillae eius super habitatores Sychem eo quod frater uester sit 9:19 si ergo recte et absque 9:19 If ye then have dealt truly and sincerely with Jerub-

baal and with his house this day, then rejoice ye in Abimelech, and let him also rejoice in you:

uitio egistis cum Hierobbaal et domo eius hodie laetamini in Abimelech et ille laetetur in uobis 9:20 sin autem per- 9:20 But if not, let fire come out from Abimelech, and

devour the men of Shechem, and the house of Millo; and let fire come out from the men of Shechem, and from the

uerse egrediatur ignis ex eo et consumat habitatores Sychem house of Millo, and devour Abimelech. et oppidum Mello egrediaturque ignis de uiris Sychem et de oppido Mello et deuoret Abimelech 9:21 quae cum dixis- 9:21 And Jotham ran away, and fled, and went to Beer, and

dwelt there, for fear of Abimelech his brother.

set fugit et abiit in Bera habitauitque ibi metu Abimelech fratris sui 9:22 regnauit itaque Abimelech super Israhel tri- 9:22 When Abimelech had reigned three years over Israel, bus annis 9:23 misitque Deus spiritum pessimum inter Abi- 9:23 Then God sent an evil spirit between Abimelech

and the men of Shechem; and the men of Shechem dealt treacherously with Abimelech:

melech et habitatores Sychem qui coeperunt eum detestari 9:24 et scelus interfectionis septuaginta filiorum Hierobbaal 9:24 That the cruelty done to the threescore and ten sons of

Jerubbaal might come, and their blood be laid upon Abimelech their brother, which slew them; and upon the men of

et effusionem sanguinis eorum conferre in Abimelech frat- Shechem, which aided him in the killing of his brethren. rem suum et in ceteros Sycimarum principes qui eum adiuuerant 9:25 posueruntque insidias aduersum eum in mon- 9:25 And the men of Shechem set liers in wait for him in

the top of the mountains, and they robbed all that came along that way by them: and it was told Abimelech.

tium summitate et dum illius praestolantur aduentum exercebant latrocinia agentes praedas de praetereuntibus nuntiatumque est Abimelech 9:26 uenit autem Gaal filius Obed 9:26 And Gaal the son of Ebed came with his brethren,

and went over to Shechem: and the men of Shechem put their confidence in him.

cum fratribus suis et transiuit in Sycimam ad cuius aduentum erecti habitatores Sychem 9:27 egressi sunt in agros 9:27 And they went out into the fields, and gathered their

vineyards, and trode the grapes, and made merry, and went into the house of their god, and did eat and drink, and

uastantes uineas uuasque calcantes et factis cantantium cho- cursed Abimelech. ris ingressi sunt fanum dei sui et inter epulas et pocula maledicebant Abimelech 9:28 clamante Gaal filio Obed quis est 9:28 And Gaal the son of Ebed said, Who is Abimelech,

and who is Shechem, that we should serve him? is not he the son of Jerubbaal? and Zebul his officer? serve the men of Hamor the father of Shechem: for why should we serve him?

Biblia Sacra Vulgata

Abimelech et quae est Sychem ut seruiamus ei numquid non est filius Hierobbaal et constituit principem Zebul seruum suum super uiros Emmor patris Sychem cur igitur seruimus

9:29 And would to God this people were under my hand! ei 9:29 utinam daret aliquis populum istum sub manu mea

then would I remove Abimelech. And he said to Abimelech, Increase thine army, and come out.

ut auferrem de medio Abimelech dictumque est Abimelech

9:30 And when Zebul the ruler of the city heard the words congrega exercitus multitudinem et ueni 9:30 Zebul enim

of Gaal the son of Ebed, his anger was kindled.

princeps ciuitatis auditis sermonibus Gaal filii Obed iratus

9:31 And he sent messengers unto Abimelech privily, say- est ualde 9:31 et misit clam ad Abimelech nuntios dicens ing, Behold, Gaal the son of Ebed and his brethren be come to Shechem; and, behold, they fortify the city against

thee. ecce Gaal filius Obed uenit in Sycimam cum fratribus suis 9:32 Now therefore up by night, thou and the people that et obpugnat aduersum te ciuitatem 9:32 surge itaque nocte

is with thee, and lie in wait in the field: 9:33 And it shall be, that in the morning, as soon as the cum populo qui tecum est et latita in agro 9:33 et primo sun is up, thou shalt rise early, and set upon the city: and, behold, when he and the people that is with him come out against thee, then mayest thou do to them as thou shalt find occasion.

mane oriente sole inrue super ciuitatem illo autem egrediente

9:34 And Abimelech rose up, and all the people that were aduersum te cum populo suo fac ei quod potueris 9:34 sur-

with him, by night, and they laid wait against Shechem in four companies.

rexit itaque Abimelech cum omni exercitu suo nocte et te-

9:35 And Gaal the son of Ebed went out, and stood in the tendit insidias iuxta Sycimam in quattuor locis 9:35 egres-

entering of the gate of the city: and Abimelech rose up, and the people that were with him, from lying in wait.

susque est Gaal filius Obed et stetit in introitu portae ciuitatis surrexit autem Abimelech et omnis exercitus cum eo de

9:36 And when Gaal saw the people, he said to Zebul, Be- insidiarum loco 9:36 cumque uidisset populum Gaal dixit

hold, there come people down from the top of the mountains. And Zebul said unto him, Thou seest the shadow of

the mountains as if they were men. ad Zebul ecce de montibus multitudo descendit cui ille res-

pondit umbras montium uides quasi hominum capita et hoc

9:37 And Gaal spake again, and said, See there come peo- errore deciperis 9:37 rursumque Gaal ait ecce populus de

ple down by the middle of the land, and another company come along by the plain of Meonenim.

umbilico terrae descendit et unus cuneus uenit per uiam quae

9:38 Then said Zebul unto him, Where is now thy mouth, respicit quercum 9:38 cui dixit Zebul ubi est nunc os tuum

wherewith thou saidst, Who is Abimelech, that we should serve him? is not this the people that thou hast despised?

go out, I pray now, and fight with them. quo loquebaris quis est Abimelech ut seruiamus ei nonne iste

est populus quem despiciebas egredere et pugna contra eum

9:39 And Gaal went out before the men of Shechem, and 9:39 abiit ergo Gaal spectante Sycimarum populo et pugna-

fought with Abimelech.

9:40 And Abimelech chased him, and he fled before him, uit contra Abimelech 9:40 qui persecutus est eum fugien-

and many were overthrown and wounded, even unto the entering of the gate.

tem et in urbem conpulit cecideruntque ex parte eius plurimi

9:41 And Abimelech dwelt at Arumah: and Zebul thrust usque ad portam ciuitatis 9:41 et Abimelech sedit in Ruma

out Gaal and his brethren, that they should not dwell in Shechem.

Zebul autem Gaal et socios eius expulit de urbe nec in ea

9:42 And it came to pass on the morrow, that the people passus est commorari 9:42 sequenti ergo die egressus est

went out into the field; and they told Abimelech.

populus in campum quod cum nuntiatum esset Abimelech

9:43 And he took the people, and divided them into three 9:43 tulit exercitum suum et diuisit in tres turmas tendens in-

companies, and laid wait in the field, and looked, and, behold, the people were come forth out of the city; and he rose up against them, and smote them.

Liber Iudicum

sidias in agris uidensque quod egrederetur populus de ciuitate surrexit et inruit in eos 9:44 cum cuneo suo obpugnans 9:44 And Abimelech, and the company that was with him,

rushed forward, and stood in the entering of the gate of the city: and the two other companies ran upon all the people

et obsidens ciuitatem duae autem turmae palantes per cam- that were in the fields, and slew them. pum aduersarios sequebantur 9:45 porro Abimelech omni 9:45 And Abimelech fought against the city all that day;

and he took the city, and slew the people that was therein, and beat down the city, and sowed it with salt.

illo die obpugnabat urbem quam cepit interfectis habitatoribus eius ipsaque destructa ita ut sal in ea dispergeret 9:46 9:46 And when all the men of the tower of Shechem heard

that, they entered into an hold of the house of the god Berith.

quod cum audissent qui habitabant in turre Sycimorum ingressi sunt fanum dei sui Berith ubi foedus cum eo pepigerant et ex eo locus nomen acceperat qui erat ualde munitus 9:47 Abimelech quoque audiens uiros turris Sycimorum pari- 9:47 And it was told Abimelech, that all the men of the

tower of Shechem were gathered together.

ter conglobatos 9:48 ascendit in montem Selmon cum omni 9:48 And Abimelech gat him up to mount Zalmon, he and

all the people that were with him; and Abimelech took an axe in his hand, and cut down a bough from the trees, and took it, and laid it on his shoulder, and said unto the people that were with him, What ye have seen me do, make haste, and do as I have done.

populo suo et arrepta securi praecidit arboris ramum inpositumque ferens umero dixit ad socios quod me uidistis facere cito facite 9:49 igitur certatim ramos de arboribus praeci- 9:49 And all the people likewise cut down every man his

bough, and followed Abimelech, and put them to the hold, and set the hold on fire upon them; so that all the men of the tower of Shechem died also, about a thousand men and women.

dentes sequebantur ducem quos circumdantes praesidio succenderunt atque ita factum est ut fumo et igne mille hominum necarentur uiri pariter ac mulieres habitatorum turris Sychem 9:50 Abimelech autem inde proficiscens uenit ad 9:50 Then went Abimelech to Thebez, and encamped

against Thebez, and took it.

oppidum Thebes quod circumdans obsidebat exercitu 9:51 9:51 But there was a strong tower within the city, and

thither fled all the men and women, and all they of the city, and shut it to them, and gat them up to the top of the

erat autem turris excelsa in media ciuitate ad quam confuge- tower. rant uiri simul ac mulieres et omnes principes ciuitatis clausa firmissime ianua et super turris tectum stantes per propugnacula 9:52 accedensque Abimelech iuxta turrem pugna- 9:52 And Abimelech came unto the tower, and fought

against it, and went hard unto the door of the tower to burn it with fire.

bat fortiter et adpropinquans ostio ignem subponere nitebatur 9:53 et ecce una mulier fragmen molae desuper iaciens 9:53 And a certain woman cast a piece of a millstone upon

Abimelech’s head, and all to brake his skull.

inlisit capiti Abimelech et confregit cerebrum eius 9:54 qui 9:54 Then he called hastily unto the young man his ar-

mourbearer, and said unto him, Draw thy sword, and slay me, that men say not of me, A women slew him. And his

uocauit cito armigerum suum et ait ad eum euagina gladium young man thrust him through, and he died. tuum et percute me ne forte dicatur quod a femina interfectus sim qui iussa perficiens interfecit eum 9:55 illoque mortuo 9:55 And when the men of Israel saw that Abimelech was

dead, they departed every man unto his place.

omnes qui cum eo erant de Israhel reuersi sunt in sedes suas 9:56 et reddidit Deus malum quod fecerat Abimelech con- 9:56 Thus God rendered the wickedness of Abimelech,

which he did unto his father, in slaying his seventy brethren:

tra patrem suum interfectis septuaginta fratribus suis 9:57

9:57 And all the evil of the men of Shechem did God render upon their heads: and upon them came the curse of Jotham the son of Jerubbaal.

Sycimitis quoque quod operati erant retributum est et uenit

Biblia Sacra Vulgata

super eos maledictio Ioatham filii Hierobbaal

10:1 And after Abimelech there arose to defend Israel Tola 10:1 post Abimelech surrexit dux in Israhel Thola filius the son of Puah, the son of Dodo, a man of Issachar; and he dwelt in Shamir in mount Ephraim.

Phoa patrui Abimelech uir de Isachar qui habitauit in Sanir

10:2 And he judged Israel twenty and three years, and montis Ephraim 10:2 et iudicauit Israhel uiginti et tribus

died, and was buried in Shamir.

10:3 And after him arose Jair, a Gileadite, and judged Is- annis mortuusque ac sepultus est in Sanir 10:3 huic succes-

rael twenty and two years.

sit Iair Galaadites qui iudicauit Israhel per uiginti et duos an-

10:4 And he had thirty sons that rode on thirty ass colts, nos 10:4 habens triginta filios sedentes super triginta pullos

and they had thirty cities, which are called Havothjair unto this day, which are in the land of Gilead.

asinarum et principes triginta ciuitatum quae ex nomine eius appellatae sunt Auothiair id est oppida Iair usque in prae-

10:5 And Jair died, and was buried in Camon. sentem diem in terra Galaad 10:5 mortuusque est Iair ac

10:6 And the children of Israel did evil again in the sight sepultus in loco cui est uocabulum Camon 10:6 filii autem of the LORD, and served Baalim, and Ashtaroth, and the gods of Syria, and the gods of Zidon, and the gods of Moab, and the gods of the children of Ammon, and the gods of the Philistines, and forsook the LORD, and served not him.

Israhel peccatis ueteribus iungentes noua fecerunt malum in conspectu Domini et seruierunt idolis Baalim et Astharoth et diis Syriae ac Sidonis et Moab et filiorum Ammon et Phi-

10:7 And the anger of the LORD was hot against Israel, listhim dimiseruntque Dominum et non colebant eum 10:7

and he sold them into the hands of the Philistines, and into the hands of the children of Ammon.

contra quos iratus tradidit eos in manu Philisthim et filio-

10:8 And that year they vexed and oppressed the children rum Ammon 10:8 adflictique sunt et uehementer oppressi of Israel: eighteen years, all the children of Israel that were on the other side Jordan in the land of the Amorites, which

is in Gilead. per annos decem et octo omnes qui habitabant trans Iorda-

10:9 Moreover the children of Ammon passed over Jor- nem in terra Amorrei quae est in Galaad 10:9 in tantum dan to fight also against Judah, and against Benjamin, and against the house of Ephraim; so that Israel was sore dis-

tressed. ut filii Ammon Iordane transmisso uastarent Iudam et Be-10:10 And the children of Israel cried unto the LORD, niamin et Ephraim adflictusque est Israhel nimis 10:10 et

saying, We have sinned against thee, both because we have forsaken our God, and also served Baalim.

clamantes ad Dominum dixerunt peccauimus tibi quia de-

10:11 And the LORD said unto the children of Israel, reliquimus Deum nostrum et seruiuimus Baalim 10:11 qui-

Did not I deliver you from the Egyptians, and from the Amorites, from the children of Ammon, and from the

Philistines? bus locutus est Dominus numquid non Aegyptii et Amor-10:12 The Zidonians also, and the Amalekites, and the rei filiique Ammon et Philisthim 10:12 Sidonii quoque et

Maonites, did oppress you; and ye cried to me, and I delivered you out of their hand.

Amalech et Chanaan oppresserunt uos et clamastis ad me et

10:13 Yet ye have forsaken me, and served other gods: erui uos de manu eorum 10:13 et tamen reliquistis me et co-

wherefore I will deliver you no more.

luistis deos alienos idcirco non addam ut ultra uos liberem

10:14 Go and cry unto the gods which ye have chosen; let 10:14 ite et inuocate deos quos elegistis ipsi uos liberent in

them deliver you in the time of your tribulation.

10:15 And the children of Israel said unto the LORD, We tempore angustiae 10:15 dixeruntque filii Israhel ad Domi-

have sinned: do thou unto us whatsoever seemeth good unto thee; deliver us only, we pray thee, this day.

num peccauimus redde tu nobis quicquid tibi placet tantum

10:16 And they put away the strange gods from among nunc libera nos 10:16 quae dicentes omnia de finibus suis

them, and served the LORD: and his soul was grieved for the misery of Israel.

alienorum deorum idola proiecerunt et seruierunt Deo qui

Liber Iudicum

doluit super miseriis eorum 10:17 itaque filii Ammon con- 10:17 Then the children of Ammon were gathered to-

gether, and encamped in Gilead. And the children of Israel assembled themselves together, and encamped in Mizpeh.

clamantes in Galaad fixere tentoria contra quos congregati filii Israhel in Maspha castrametati sunt 10:18 dixeruntque 10:18 And the people and princes of Gilead said one to

another, What man is he that will begin to fight against the children of Ammon? he shall be head over all the inhabi-

principes Galaad singuli ad proximos suos qui primus e no- tants of Gilead. bis contra filios Ammon coeperit dimicare erit dux populi Galaad

11:1 fuit illo tempore Iepthae Galaadites uir fortissimus 11:1 Now Jephthah the Gileadite was a mighty man of val-

our, and he was the son of an harlot: and Gilead begat Jephthah.

atque pugnator filius meretricis mulieris qui natus est de Ga-

laad 11:2 habuit autem Galaad uxorem de qua suscepit fi- 11:2 And Gilead’s wife bare him sons; and his wife’s sons grew up, and they thrust out Jephthah, and said unto him, Thou shalt not inherit in our father’s house; for thou art

lios qui postquam creuerant eiecerunt Iepthae dicentes heres the son of a strange woman. in domo patris nostri esse non poteris quia de altera matre generatus es 11:3 quos ille fugiens atque deuitans habitauit 11:3 Then Jephthah fled from his brethren, and dwelt in

the land of Tob: and there were gathered vain men to Jephthah, and went out with him.

in terra Tob congregatique sunt ad eum uiri inopes et latrocinantes et quasi principem sequebantur 11:4 in illis diebus 11:4 And it came to pass in process of time, that the chil-

dren of Ammon made war against Israel.

pugnabant filii Ammon contra Israhel 11:5 quibus acriter 11:5 And it was so, that when the children of Ammon

made war against Israel, the elders of Gilead went to fetch Jephthah out of the land of Tob:

instantibus perrexerunt maiores natu de Galaad ut tollerent in auxilium sui Iepthae de terra Tob 11:6 dixeruntque ad 11:6 And they said unto Jephthah, Come, and be our cap-

tain, that we may fight with the children of Ammon.

eum ueni et esto princeps noster et pugna contra filios Ammon 11:7 quibus ille respondit nonne uos estis qui odistis 11:7 And Jephthah said unto the elders of Gilead, Did not

ye hate me, and expel me out of my father’s house? and why are ye come unto me now when ye are in distress?

me et eiecistis de domo patris mei et nunc uenistis ad me necessitate conpulsi 11:8 dixeruntque principes Galaad ad 11:8 And the elders of Gilead said unto Jephthah, There-

fore we turn again to thee now, that thou mayest go with us, and fight against the children of Ammon, and be our

Iepthae ob hanc igitur causam nunc ad te uenimus ut profici- head over all the inhabitants of Gilead. scaris nobiscum et pugnes contra filios Ammon sisque dux omnium qui habitant in Galaad 11:9 Iepthae quoque dixit 11:9 And Jephthah said unto the elders of Gilead, If ye

bring me home again to fight against the children of Ammon, and the LORD deliver them before me, shall I be

eis si uere uenistis ad me ut pugnem pro uobis contra fi- your head? lios Ammon tradideritque eos Dominus in manus meas ego ero princeps uester 11:10 qui responderunt ei Dominus qui 11:10 And the elders of Gilead said unto Jephthah, The

LORD be witness between us, if we do not so according to thy words.

haec audit ipse mediator ac testis est quod nostra promissa faciamus 11:11 abiit itaque Iepthae cum principibus Ga- 11:11 Then Jephthah went with the elders of Gilead,

and the people made him head and captain over them: and Jephthah uttered all his words before the LORD in

laad fecitque eum omnis populus principem sui locutusque Mizpeh. est Iepthae omnes sermones suos coram Domino in Maspha 11:12 et misit nuntios ad regem filiorum Ammon qui ex per- 11:12 And Jephthah sent messengers unto the king of the

children of Ammon, saying, What hast thou to do with me, that thou art come against me to fight in my land?

sona sua dicerent quid mihi et tibi est quia uenisti contra me

Biblia Sacra Vulgata

11:13 And the king of the children of Ammon answered ut uastares terram meam 11:13 quibus ille respondit quia

unto the messengers of Jephthah, Because Israel took away my land, when they came up out of Egypt, from Arnon even unto Jabbok, and unto Jordan: now therefore restore those lands again peaceably.

tulit Israhel terram meam quando ascendit de Aegypto a finibus Arnon usque Iaboc atque Iordanem nunc igitur cum pace

11:14 And Jephthah sent messengers again unto the king redde mihi eam 11:14 per quos rursum mandauit Iepthae et

of the children of Ammon:

11:15 And said unto him, Thus saith Jephthah, Israel took imperauit eis ut dicerent regi Ammon 11:15 haec dicit Iep-

not away the land of Moab, nor the land of the children of Ammon:

thae non tulit Israhel terram Moab nec terram filiorum Am-

11:16 But when Israel came up from Egypt, and walked mon 11:16 sed quando de Aegypto conscenderunt ambula-

through the wilderness unto the Red sea, and came to Kadesh;

uit per solitudinem usque ad mare Rubrum et uenit in Ca-

11:17 Then Israel sent messengers unto the king of Edom, des 11:17 misitque nuntios ad regem Edom dicens dimitte saying, Let me, I pray thee, pass through thy land: but the king of Edom would not hearken thereto. And in like manner they sent unto the king of Moab: but he would not consent: and Israel abode in Kadesh.

ut transeam per terram tuam qui noluit adquiescere precibus eius misit quoque et ad regem Moab qui et ipse trans-

11:18 Then they went along through the wilderness, and itum praebere contempsit mansit itaque in Cades 11:18 et

compassed the land of Edom, and the land of Moab, and came by the east side of the land of Moab, and pitched on the other side of Arnon, but came not within the border of Moab: for Arnon was the border of Moab.

circuiuit ex latere terram Edom et terram Moab uenitque contra orientalem plagam terrae Moab et castrametatus est trans Arnon nec uoluit intrare terminos Moab Arnon quippe

11:19 And Israel sent messengers unto Sihon king of the confinium est terrae Moab 11:19 misit itaque Israhel nun-

Amorites, the king of Heshbon; and Israel said unto him, Let us pass, we pray thee, through thy land into my place.

tios ad Seon regem Amorreorum qui habitabat in Esebon et dixerunt ei dimitte ut transeam per terram tuam usque ad flu-

11:20 But Sihon trusted not Israel to pass through his uium 11:20 qui et ipse Israhel uerba despiciens non dimisit

coast: but Sihon gathered all his people together, and pitched in Jahaz, and fought against Israel.

eum transire per terminos suos sed infinita multitudine congregata egressus est contra eum in Iassa et fortiter resistebat

11:21 And the LORD God of Israel delivered Sihon and 11:21 tradiditque eum Dominus in manu Israhel cum omni

all his people into the hand of Israel, and they smote them: so Israel possessed all the land of the Amorites, the inhab-

itants of that country. exercitu suo qui percussit eum et possedit omnem terram 11:22 And they possessed all the coasts of the Amorites, Amorrei habitatoris regionis illius 11:22 et uniuersos fines

from Arnon even unto Jabbok, and from the wilderness even unto Jordan.

eius de Arnon usque Iaboc et de solitudine usque ad Iorda-

11:23 So now the LORD God of Israel hath dispossessed nem 11:23 Dominus ergo Deus Israhel subuertit Amorreum

the Amorites from before his people Israel, and shouldest thou possess it?

pugnante contra illum populo suo Israhel et tu nunc uis pos-

11:24 Wilt not thou possess that which Chemosh thy god sidere terram eius 11:24 nonne ea quae possedit Chamos

giveth thee to possess? So whomsoever the LORD our God shall drive out from before us, them will we possess.

deus tuus tibi iure debentur quae autem Dominus Deus no-

11:25 And now art thou any thing better than Balak the ster uictor obtinuit in nostram cedent possessionem 11:25

son of Zippor, king of Moab? did he ever strive against Israel, or did he ever fight against them,

nisi forte melior es Balac filio Sepphor rege Moab aut docere potes quod iurgatus sit contra Israhel et pugnauerit contra

11:26 While Israel dwelt in Heshbon and her towns, and in eum 11:26 quando habitauit in Esebon et uiculis eius et in

Aroer and her towns, and in all the cities that be along by the coasts of Arnon, three hundred years? why therefore did ye not recover them within that time?

Liber Iudicum

Aroer et uillis illius uel in cunctis ciuitatibus iuxta Iordanem per trecentos annos quare tanto tempore nihil super hac repetitione temptastis 11:27 igitur non ego pecco in te sed tu 11:27 Wherefore I have not sinned against thee, but thou

doest me wrong to war against me: the LORD the Judge be judge this day between the children of Israel and the

contra me male agis indicens mihi bella non iusta iudicet Do- children of Ammon. minus arbiter huius diei inter Israhel et inter filios Ammon 11:28 noluitque adquiescere rex filiorum Ammon uerbis Iep- 11:28 Howbeit the king of the children of Ammon hear-

kened not unto the words of Jephthah which he sent him.

thae quae per nuntios mandauerat 11:29 factus est ergo su- 11:29 Then the Spirit of the LORD came upon Jephthah,

and he passed over Gilead, and Manasseh, and passed over Mizpeh of Gilead, and from Mizpeh of Gilead he passed

per Iepthae spiritus Domini et circumiens Galaad et Manasse over unto the children of Ammon. Maspha quoque Galaad et inde transiens ad filios Ammon 11:30 uotum uouit Domino dicens si tradideris filios Ammon 11:30 And Jephthah vowed a vow unto the LORD, and

said, If thou shalt without fail deliver the children of Ammon into mine hands,

in manus meas 11:31 quicumque primus fuerit egressus de

11:31 Then it shall be, that whatsoever cometh forth of the doors of my house to meet me, when I return in peace from the children of Ammon, shall surely be the LORD’s, and I will offer it up for a burnt offering.

foribus domus meae mihique occurrerit reuertenti cum pace

a filiis Ammon eum holocaustum offeram Domino 11:32 11:32 So Jephthah passed over unto the children of Ammon to fight against them; and the LORD delivered them into his hands.

transiuitque Iepthae ad filios Ammon ut pugnaret contra eos quos tradidit Dominus in manus eius 11:33 percussitque ab 11:33 And he smote them from Aroer, even till thou come

to Minnith, even twenty cities, and unto the plain of the vineyards, with a very great slaughter. Thus the children

Aroer usque dum uenias in Mennith uiginti ciuitates et usque of Ammon were subdued before the children of Israel. ad Abel quae est uineis consita plaga magna nimis humiliatique sunt filii Ammon a filiis Israhel 11:34 reuertenti autem 11:34 And Jephthah came to Mizpeh unto his house, and,

behold, his daughter came out to meet him with timbrels and with dances: and she was his only child; beside her he

Iepthae in Maspha domum suam occurrit unigenita filia cum had neither son nor daughter. tympanis et choris non enim habebat alios liberos 11:35 qua 11:35 And it came to pass, when he saw her, that he rent

his clothes, and said, Alas, my daughter! thou hast brought me very low, and thou art one of them that trouble me: for I have opened my mouth unto the LORD, and I cannot go back.

uisa scidit uestimenta sua et ait heu filia mi decepisti me et ipsa decepta es aperui enim os meum ad Dominum et aliud facere non potero 11:36 cui illa respondit pater mi si 11:36 And she said unto him, My father, if thou hast

opened thy mouth unto the LORD, do to me according to that which hath proceeded out of thy mouth; forasmuch as the LORD hath taken vengeance for thee of thine enemies, even of the children of Ammon.

aperuisti os tuum ad Dominum fac mihi quodcumque pollicitus es concessa tibi ultione atque uictoria de hostibus tuis 11:37 dixitque ad patrem hoc solum mihi praesta quod de- 11:37 And she said unto her father, Let this thing be done

for me: let me alone two months, that I may go up and down upon the mountains, and bewail my virginity, I and

precor dimitte me ut duobus mensibus circumeam montes et my fellows. plangam uirginitatem meam cum sodalibus meis 11:38 cui 11:38 And he said, Go. And he sent her away for two

months: and she went with her companions, and bewailed her virginity upon the mountains.

ille respondit uade et dimisit eam duobus mensibus cumque abisset cum sociis ac sodalibus suis flebat uirginitatem suam in montibus 11:39 expletisque duobus mensibus reuersa est 11:39 And it came to pass at the end of two months, that

she returned unto her father, who did with her according to his vow which he had vowed: and she knew no man.

ad patrem suum et fecit ei sicut uouerat quae ignorabat ui- And it was a custom in Israel, rum exinde mos increbuit in Israhel et consuetudo seruata

Biblia Sacra Vulgata

11:40 That the daughters of Israel went yearly to lament est 11:40 ut post anni circulum conueniant in unum filiae the daughter of Jephthah the Gileadite four days in a year.

Israhel et plangant filiam Iepthae Galaaditae diebus quattuor

12:1 And the men of Ephraim gathered themselves to- 12:1 ecce autem in Ephraim orta seditio est nam transeun-

gether, and went northward, and said unto Jephthah, Wherefore passedst thou over to fight against the children of Ammon, and didst not call us to go with thee? we will burn thine house upon thee with fire.

tes contra aquilonem dixerunt ad Iepthae quare uadens ad pugnam contra filios Ammon uocare nos noluisti ut pergere-

12:2 And Jephthah said unto them, I and my people were mus tecum igitur incendimus domum tuam 12:2 quibus ille

at great strife with the children of Ammon; and when I called you, ye delivered me not out of their hands.

respondit disceptatio erat mihi et populo meo contra filios Ammon uehemens uocauique uos ut mihi praeberetis auxi-

12:3 And when I saw that ye delivered me not, I put my life lium et facere noluistis 12:3 quod cernens posui in manibus

in my hands, and passed over against the children of Ammon, and the LORD delivered them into my hand: wherefore then are ye come up unto me this day, to fight against me?

meis animam meam transiuique ad filios Ammon et tradidit eos Dominus in manus meas quid commerui ut aduer-

12:4 Then Jephthah gathered together all the men of sum me consurgatis in proelium 12:4 uocatis itaque ad se

Gilead, and fought with Ephraim: and the men of Gilead smote Ephraim, because they said, Ye Gileadites are fugitives of Ephraim among the Ephraimites, and among the Manassites.

cunctis uiris Galaad pugnabat contra Ephraim percusseruntque uiri Galaad Ephraim quia dixerat fugitiuus est Galaad de

12:5 And the Gileadites took the passages of Jordan be- Ephraim et habitat in medio Ephraim et Manasse 12:5 ocfore the Ephraimites: and it was so, that when those Ephraimites which were escaped said, Let me go over; that the men of Gilead said unto him, Art thou an Ephraimite? If he said, Nay;

cupaueruntque Galaaditae uada Iordanis per quae Ephraim reuersurus erat cumque uenisset ad ea de Ephraim numero fugiens atque dixisset obsecro ut me transire permittas dicebant ei Galaaditae numquid Ephrateus es quo dicente non

12:6 Then said they unto him, Say now Shibboleth: and sum 12:6 interrogabant eum dic ergo sebboleth quod inter-

he said Sibboleth: for he could not frame to pronounce it right. Then they took him, and slew him at the passages of Jordan: and there fell at that time of the Ephraimites forty and two thousand.

pretatur spica qui respondebat tebboleth eadem littera spicam exprimere non ualens statimque adprehensum iugulabant in ipso Iordanis transitu et ceciderunt in illo tempore

12:7 And Jephthah judged Israel six years. Then died de Ephraim quadraginta duo milia 12:7 iudicauitque Iep-

Jephthah the Gileadite, and was buried in one of the cities of Gilead.

thae Galaadites Israhel sex annis et mortuus est ac sepul-

12:8 And after him Ibzan of Bethlehem judged Israel. tus in ciuitate sua Galaad 12:8 post hunc iudicauit Israhel 12:9 And he had thirty sons, and thirty daughters, whom Abessan de Bethleem 12:9 qui habuit triginta filios et toti-

he sent abroad, and took in thirty daughters from abroad for his sons. And he judged Israel seven years.

dem filias quas emittens foras maritis dedit et eiusdem numeri filiis suis accepit uxores introducens in domum suam

12:10 Then died Ibzan, and was buried at Bethlehem. qui septem annis iudicauit Israhel 12:10 mortuusque est ac

12:11 And after him Elon, a Zebulonite, judged Israel; and sepultus in Bethleem 12:11 cui successit Ahialon Zabulohe judged Israel ten years.

12:12 And Elon the Zebulonite died, and was buried in nites et iudicauit Israhelem decem annis 12:12 mortuusque

Aijalon in the country of Zebulun.

12:13 And after him Abdon the son of Hillel, a Pirathonite, est ac sepultus in Zabulon 12:13 post hunc iudicauit in Isra-

judged Israel.

Liber Iudicum

hel Abdon filius Hellel Farathonites 12:14 qui habuit quad- 12:14 And he had forty sons and thirty nephews, that rode

on threescore and ten ass colts: and he judged Israel eight years.

raginta filios et triginta ex eis nepotes ascendentes super septuaginta pullos asinarum et iudicauit in Israhel octo annis 12:15 mortuusque est ac sepultus in Farathon terrae Ephraim 12:15 And Abdon the son of Hillel the Pirathonite died,

and was buried in Pirathon in the land of Ephraim, in the mount of the Amalekites.

in monte Amalech

13:1 rursumque filii Israhel fecerunt malum in conspectu 13:1 And the children of Israel did evil again in the sight

of the LORD; and the LORD delivered them into the hand of the Philistines forty years.

Domini qui tradidit eos in manus Philisthinorum quadraginta annis 13:2 erat autem uir quidam de Saraa et de stirpe Dan 13:2 And there was a certain man of Zorah, of the family

of the Danites, whose name was Manoah; and his wife was barren, and bare not.

nomine Manue habens uxorem sterilem 13:3 cui apparuit

13:3 And the angel of the LORD appeared unto the woman, and said unto her, Behold now, thou art barren, and bearest not: but thou shalt conceive, and bear a son.

angelus Domini et dixit ad eam sterilis es et absque liberis sed concipies et paries filium 13:4 caue ergo ne uinum bi- 13:4 Now therefore beware, I pray thee, and drink not

wine nor strong drink, and eat not any unclean thing:

bas ac siceram ne inmundum quicquam comedas 13:5 quia 13:5 For, lo, thou shalt conceive, and bear a son; and no razor shall come on his head: for the child shall be a Nazarite unto God from the womb: and he shall begin to deliver Is-

concipies et paries filium cuius non tanget caput nouacula rael out of the hand of the Philistines. erit enim nazareus Dei ab infantia sua et ex matris utero et ipse incipiet liberare Israhel de manu Philisthinorum 13:6 13:6 Then the woman came and told her husband, saying,

A man of God came unto me, and his countenance was like the countenance of an angel of God, very terrible: but I asked him not whence he was, neither told he me his name:

quae cum uenisset ad maritum dixit ei uir Dei uenit ad me habens uultum angelicum terribilis nimis quem cum interrogassem quis esset et unde uenisset et quo nomine uocaretur noluit mihi dicere 13:7 sed hoc respondit ecce concipies et 13:7 But he said unto me, Behold, thou shalt conceive, and

bear a son; and now drink no wine nor strong drink, neither eat any unclean thing: for the child shall be a Nazarite to

paries filium caue ne uinum bibas et siceram et ne aliquo God from the womb to the day of his death. uescaris inmundo erit enim puer nazareus Dei ab infantia sua et ex utero matris usque ad diem mortis suae 13:8 ora- 13:8 Then Manoah intreated the LORD, and said, O my

Lord, let the man of God which thou didst send come again unto us, and teach us what we shall do unto the child that

uit itaque Manue Deum et ait obsecro Domine ut uir Dei shall be born. quem misisti ueniat iterum et doceat nos quid debeamus facere de puero qui nasciturus est 13:9 exaudiuitque Dominus 13:9 And God hearkened to the voice of Manoah; and the

angel of God came again unto the woman as she sat in the field: but Manoah her husband was not with her.

precantem Manue et apparuit rursum angelus Domini uxori eius sedenti in agro Manue autem maritus eius non erat cum ea quae cum uidisset angelum 13:10 festinauit et cucurrit 13:10 And the woman made haste, and ran, and shewed

her husband, and said unto him, Behold, the man hath appeared unto me, that came unto me the other day.

ad uirum suum nuntiauitque ei dicens ecce apparuit mihi uir quem ante uideram 13:11 qui surrexit et secutus est uxo- 13:11 And Manoah arose, and went after his wife, and

came to the man, and said unto him, Art thou the man that spakest unto the woman? And he said, I am.

rem suam ueniensque ad uirum dixit ei tu es qui locutus es mulieri et ille respondit ego sum 13:12 cui Manue quando 13:12 And Manoah said, Now let thy words come to pass.

How shall we order the child, and how shall we do unto him?

inquit sermo tuus fuerit expletus quid uis ut faciat puer aut a

Biblia Sacra Vulgata

13:13 And the angel of the LORD said unto Manoah, Of quo se obseruare debebit 13:13 dixitque angelus Domini ad

all that I said unto the woman let her beware.

Manue ab omnibus quae locutus sum uxori tuae abstineat se

13:14 She may not eat of any thing that cometh of the 13:14 et quicquid ex uinea nascitur non comedat uinum et si-

vine, neither let her drink wine or strong drink, nor eat any unclean thing: all that I commanded her let her observe.

ceram non bibat nullo uescatur inmundo et quod ei praecepi

13:15 And Manoah said unto the angel of the LORD, I impleat atque custodiat 13:15 dixitque Manue ad angelum

pray thee, let us detain thee, until we shall have made ready a kid for thee.

Domini obsecro te ut adquiescas precibus meis et faciamus

13:16 And the angel of the LORD said unto Manoah, tibi hedum de capris 13:16 cui respondit angelus si me co-

Though thou detain me, I will not eat of thy bread: and if thou wilt offer a burnt offering, thou must offer it unto the LORD. For Manoah knew not that he was an angel of the LORD.

gis non comedam panes tuos sin autem uis holocaustum facere offer illud Domino et nesciebat Manue quod angelus

13:17 And Manoah said unto the angel of the LORD, Dei esset 13:17 dixitque ad eum quod est tibi nomen ut si

What is thy name, that when thy sayings come to pass we may do thee honour?

sermo tuus fuerit expletus honoremus te 13:18 cui ille re-

13:18 And the angel of the LORD said unto him, Why askest thou thus after my name, seeing it is secret?

13:19 So Manoah took a kid with a meat offering, and spondit cur quaeris nomen meum quod est mirabile 13:19

offered it upon a rock unto the LORD: and the angel did wonderously; and Manoah and his wife looked on.

tulit itaque Manue hedum de capris et libamenta et posuit super petram offerens Domino qui facit mirabilia ipse autem

13:20 For it came to pass, when the flame went up toward et uxor eius intuebantur 13:20 cumque ascenderet flamma

heaven from off the altar, that the angel of the LORD ascended in the flame of the altar. And Manoah and his wife

looked on it, and fell on their faces to the ground. altaris in caelum angelus Domini in flamma pariter ascen-

dit quod cum uidisset Manue et uxor eius proni ceciderunt

13:21 But the angel of the LORD did no more appear to in terram 13:21 et ultra non eis apparuit angelus Domini

Manoah and to his wife. Then Manoah knew that he was an angel of the LORD.

statimque intellexit Manue angelum esse Domini 13:22 et

13:22 And Manoah said unto his wife, We shall surely die, because we have seen God.

dixit ad uxorem suam morte moriemur quia uidimus Deum

13:23 But his wife said unto him, If the LORD were 13:23 cui respondit mulier si Dominus nos uellet occidere de

pleased to kill us, he would not have received a burnt offering and a meat offering at our hands, neither would he have shewed us all these things, nor would as at this time have told us such things as these.

manibus nostris holocaustum et libamenta non suscepisset nec ostendisset nobis haec omnia neque ea quae sunt uen-

13:24 And the woman bare a son, and called his name tura dixisset 13:24 peperit itaque filium et uocauit nomen

Samson: and the child grew, and the LORD blessed him.

13:25 And the Spirit of the LORD began to move him at eius Samson creuitque puer et benedixit ei Dominus 13:25

times in the camp of Dan between Zorah and Eshtaol.

coepitque spiritus Domini esse cum eo in castris Dan inter Saraa et Esthaol

14:1 And Samson went down to Timnath, and saw a 14:1 descendit igitur Samson in Thamnatha uidensque ibi

woman in Timnath of the daughters of the Philistines.

14:2 And he came up, and told his father and his mother, mulierem de filiabus Philisthim 14:2 ascendit et nuntiauit

and said, I have seen a woman in Timnath of the daughters of the Philistines: now therefore get her for me to wife.

patri suo et matri dicens uidi mulierem in Thamnatha de filiabus Philisthinorum quam quaeso ut mihi accipiatis uxo-

14:3 Then his father and his mother said unto him, Is there rem 14:3 cui dixerunt pater et mater sua numquid non est never a woman among the daughters of thy brethren, or among all my people, that thou goest to take a wife of the uncircumcised Philistines? And Samson said unto his father, Get her for me; for she pleaseth me well.

mulier in filiabus fratrum tuorum et in omni populo meo

Liber Iudicum

quia uis accipere uxorem de Philisthim qui incircumcisi sunt dixitque Samson ad patrem suum hanc mihi accipe quia placuit oculis meis 14:4 parentes autem eius nesciebant quod 14:4 But his father and his mother knew not that it was

of the LORD, that he sought an occasion against the Philistines: for at that time the Philistines had dominion

res a Domino fieret et quaereret occasionem contra Phi- over Israel. listhim eo enim tempore Philisthim dominabantur Israheli 14:5 descendit itaque Samson cum patre suo et matre in 14:5 Then went Samson down, and his father and his

mother, to Timnath, and came to the vineyards of Timnath: and, behold, a young lion roared against him.

Thamnatha cumque uenissent ad uineas oppidi apparuit catulus leonis saeuus rugiens et occurrit ei 14:6 inruit autem 14:6 And the Spirit of the LORD came mightily upon him,

and he rent him as he would have rent a kid, and he had nothing in his hand: but he told not his father or his mother

spiritus Domini in Samson et dilacerauit leonem quasi he- what he had done. dum in frusta concerperet nihil omnino habens in manu et hoc patri et matri noluit indicare 14:7 descenditque et locu- 14:7 And he went down, and talked with the woman; and

she pleased Samson well.

tus est mulieri quae placuerat oculis eius 14:8 et post ali- 14:8 And after a time he returned to take her, and he turned

aside to see the carcase of the lion: and, behold, there was a swarm of bees and honey in the carcase of the lion.

quot dies reuertens ut acciperet eam declinauit ut uideret cadauer leonis et ecce examen apium in ore leonis erat ac fauus mellis 14:9 quem cum sumpsisset in manibus comedebat in 14:9 And he took thereof in his hands, and went on eating,

and came to his father and mother, and he gave them, and they did eat: but he told not them that he had taken the

uia ueniensque ad patrem suum et matrem dedit eis partem honey out of the carcase of the lion. qui et ipsi comederunt nec tamen eis uoluit indicare quod mel de corpore leonis adsumpserat 14:10 descendit itaque 14:10 So his father went down unto the woman: and Sam-

son made there a feast; for so used the young men to do.

pater eius ad mulierem et fecit filio suo Samson conuiuium sic enim iuuenes facere consuerant 14:11 cum igitur ciues 14:11 And it came to pass, when they saw him, that they

brought thirty companions to be with him.

loci uidissent eum dederunt ei sodales triginta qui essent cum eo 14:12 quibus locutus est Samson proponam uobis 14:12 And Samson said unto them, I will now put forth a

riddle unto you: if ye can certainly declare it me within the seven days of the feast, and find it out, then I will give

problema quod si solueritis mihi intra septem dies conuiuii you thirty sheets and thirty change of garments: dabo uobis triginta sindones et totidem tunicas 14:13 sin 14:13 But if ye cannot declare it me, then shall ye give me

thirty sheets and thirty change of garments. And they said unto him, Put forth thy riddle, that we may hear it.

autem non potueritis soluere uos dabitis mihi triginta sindones et eiusdem numeri tunicas qui responderunt ei propone problema ut audiamus 14:14 dixitque eis de comedente ex- 14:14 And he said unto them, Out of the eater came forth

meat, and out of the strong came forth sweetness. And they could not in three days expound the riddle.

iuit cibus et de forte est egressa dulcedo nec potuerunt per tres dies propositionem soluere 14:15 cumque adesset dies 14:15 And it came to pass on the seventh day, that they

said unto Samson’s wife, Entice thy husband, that he may declare unto us the riddle, lest we burn thee and thy father’s house with fire: have ye called us to take that we have? is it not so?

septimus dixerunt ad uxorem Samson blandire uiro tuo et suade ei ut indicet tibi quid significet problema quod si facere nolueris incendimus et te et domum patris tui an idcirco nos uocastis ad nuptias ut spoliaretis 14:16 quae fundebat 14:16 And Samson’s wife wept before him, and said, Thou

dost but hate me, and lovest me not: thou hast put forth a riddle unto the children of my people, and hast not told it me. And he said unto her, Behold, I have not told it my father nor my mother, and shall I tell it thee?

Biblia Sacra Vulgata

apud Samson lacrimas et querebatur dicens odisti me et non diligis idcirco problema quod proposuisti filiis populi mei non uis mihi exponere at ille respondit patri meo et matri no-

14:17 And she wept before him the seven days, while their lui dicere et tibi indicare potero 14:17 septem igitur diebus

feast lasted: and it came to pass on the seventh day, that he told her, because she lay sore upon him: and she told

the riddle to the children of her people. conuiuii flebat apud eum tandemque die septimo cum ei mo-14:18 And the men of the city said unto him on the sev- lesta esset exposuit quae statim indicauit ciuibus suis 14:18

enth day before the sun went down, What is sweeter than honey? And what is stronger than a lion? and he said unto them, If ye had not plowed with my heifer, ye had not found out my riddle.

et illi dixerunt ei die septimo ante solis occubitum quid dulcius melle et quid leone fortius qui ait ad eos si non arassetis

14:19 And the Spirit of the LORD came upon him, and he in uitula mea non inuenissetis propositionem meam 14:19

went down to Ashkelon, and slew thirty men of them, and took their spoil, and gave change of garments unto them which expounded the riddle. And his anger was kindled, and he went up to his father’s house.

inruit itaque in eo spiritus Domini descenditque Ascalonem et percussit ibi triginta uiros quorum ablatas uestes dedit his qui problema soluerant iratusque nimis ascendit in domum

14:20 But Samson’s wife was given to his companion, patris sui 14:20 uxor autem eius accepit maritum unum de

whom he had used as his friend.

amicis eius et pronubis

15:1 But it came to pass within a while after, in the time 15:1 post aliquantum autem temporis cum dies triticeae

of wheat harvest, that Samson visited his wife with a kid; and he said, I will go in to my wife into the chamber. But

her father would not suffer him to go in. messis instarent uenit Samson inuisere uolens uxorem suam

et adtulit ei hedum de capris cumque cubiculum eius solito

15:2 And her father said, I verily thought that thou hadst uellet intrare prohibuit eum pater illius dicens 15:2 putaui

utterly hated her; therefore I gave her to thy companion: is not her younger sister fairer than she? take her, I pray

thee, instead of her. quod odisses eam et ideo tradidi illam amico tuo sed habet

sororem quae iunior et pulchrior illa est sit tibi pro ea uxor

15:3 And Samson said concerning them, Now shall I be 15:3 cui respondit Samson ab hac die non erit culpa in me

more blameless than the Philistines, though I do them a displeasure.

contra Philistheos faciam enim uobis mala 15:4 perrexitque

15:4 And Samson went and caught three hundred foxes, and took firebrands, and turned tail to tail, and put a firebrand in the midst between two tails.

et cepit trecentas uulpes caudasque earum iunxit ad caudas

15:5 And when he had set the brands on fire, he let them go et faces ligauit in medio 15:5 quas igne succendens dimisit

into the standing corn of the Philistines, and burnt up both the shocks, and also the standing corn, with the vineyards

and olives. ut huc illucque discurrerent quae statim perrexerunt in segetes Philisthinorum quibus succensis et conportatae iam fruges et adhuc stantes in stipula concrematae sunt in tantum ut

15:6 Then the Philistines said, Who hath done this? And uineas quoque et oliueta flamma consumeret 15:6 dixerunt-

they answered, Samson, the son in law of the Timnite, because he had taken his wife, and given her to his companion. And the Philistines came up, and burnt her and her father with fire.

que Philisthim quis fecit hanc rem quibus dictum est Samson gener Thamnathei quia tulit uxorem eius et alteri tradidit haec operatus est ascenderuntque Philisthim et conbuserunt

15:7 And Samson said unto them, Though ye have done tam mulierem quam patrem eius 15:7 quibus ait Samson li-

this, yet will I be avenged of you, and after that I will cease.

cet haec feceritis tamen adhuc ex uobis expetam ultionem

Liber Iudicum

et tunc quiescam 15:8 percussitque eos ingenti plaga ita ut 15:8 And he smote them hip and thigh with a great slaugh-

ter: and he went down and dwelt in the top of the rock Etam.

stupentes suram femori inponerent et descendens habitauit in spelunca petrae Aetham 15:9 igitur ascendentes Philist- 15:9 Then the Philistines went up, and pitched in Judah,

and spread themselves in Lehi.

him in terra Iuda castrametati sunt et in loco qui postea uocatus est Lehi id est Maxilla eorum est fusus exercitus 15:10 15:10 And the men of Judah said, Why are ye come up

against us? And they answered, To bind Samson are we come up, to do to him as he hath done to us.

dixeruntque ad eos de tribu Iuda cur ascendistis aduersum nos qui responderunt ut ligemus Samson uenimus et reddamus ei quae in nos operatus est 15:11 descenderunt ergo 15:11 Then three thousand men of Judah went to the top

of the rock Etam, and said to Samson, Knowest thou not that the Philistines are rulers over us? what is this that thou hast done unto us? And he said unto them, As they did unto me, so have I done unto them.

tria milia uirorum de Iuda ad specum silicis Aetham dixeruntque ad Samson nescis quod Philisthim imperent nobis quare hoc facere uoluisti quibus ille ait sicut fecerunt mihi feci eis 15:12 ligare inquiunt te uenimus et tradere in ma- 15:12 And they said unto him, We are come down to

bind thee, that we may deliver thee into the hand of the Philistines. And Samson said unto them, Swear unto me,

nus Philisthinorum iurate respondit mihi quod non me oc- that ye will not fall upon me yourselves. cidatis 15:13 dixerunt non te occidimus sed uinctum tradi- 15:13 And they spake unto him, saying, No; but we will

bind thee fast, and deliver thee into their hand: but surely we will not kill thee. And they bound him with two new

mus ligaueruntque eum duobus nouis funibus et tulerunt de cords, and brought him up from the rock. petra Aetham 15:14 qui cum uenisset ad locum Maxillae 15:14 And when he came unto Lehi, the Philistines

shouted against him: and the Spirit of the LORD came mightily upon him, and the cords that were upon his arms became as flax that was burnt with fire, and his bands loosed from off his hands.

et Philisthim uociferantes occurrissent ei inruit spiritus Domini in eum et sicut solent ad odorem ignis lina consumi ita uincula quibus ligatus erat dissipata sunt et soluta 15:15 in- 15:15 And he found a new jawbone of an ass, and put forth

his hand, and took it, and slew a thousand men therewith.

uentamque maxillam id est mandibulam asini quae iacebat arripiens interfecit in ea mille uiros 15:16 et ait in maxilla 15:16 And Samson said, With the jawbone of an ass, heaps

upon heaps, with the jaw of an ass have I slain a thousand men.

asini in mandibula pulli asinarum deleui eos et percussi mille uiros 15:17 cumque haec canens uerba conplesset proiecit 15:17 And it came to pass, when he had made an end of

speaking, that he cast away the jawbone out of his hand, and called that place Ramathlehi.

mandibulam de manu et uocauit nomen loci illius Ramathlehi quod interpretatur eleuatio Maxillae 15:18 sitiensque 15:18 And he was sore athirst, and called on the LORD,

and said, Thou hast given this great deliverance into the hand of thy servant: and now shall I die for thirst, and fall

ualde clamauit ad Dominum et ait tu dedisti in manu serui into the hand of the uncircumcised? tui salutem hanc maximam atque uictoriam et en siti morior incidamque in manus incircumcisorum 15:19 aperuit itaque 15:19 But God clave an hollow place that was in the jaw,

and there came water thereout; and when he had drunk, his spirit came again, and he revived: wherefore he called the

Dominus molarem dentem in maxilla asini et egressae sunt name thereof Enhakkore, which is in Lehi unto this day. ex eo aquae quibus haustis refocilauit spiritum et uires recepit idcirco appellatum est nomen loci illius Fons inuocantis de maxilla usque in praesentem diem 15:20 iudicauitque Is- 15:20 And he judged Israel in the days of the Philistines

twenty years.

rahel in diebus Philisthim uiginti annis

Biblia Sacra Vulgata

16:1 Then went Samson to Gaza, and saw there an harlot, 16:1 abiit quoque in Gazam et uidit ibi meretricem muand went in unto her. 16:2 And it was told the Gazites, saying, Samson is come lierem ingressusque est ad eam 16:2 quod cum audissent hither. And they compassed him in, and laid wait for him all night in the gate of the city, and were quiet all the night,

saying, In the morning, when it is day, we shall kill him. Philisthim et percrebruisset apud eos intrasse urbem Samson circumdederunt eum positis in porta ciuitatis custodibus et ibi tota nocte cum silentio praestolantes ut facto mane ex-

16:3 And Samson lay till midnight, and arose at midnight, euntem occiderent 16:3 dormiuit autem Samson usque ad and took the doors of the gate of the city, and the two posts, and went away with them, bar and all, and put them upon his shoulders, and carried them up to the top of an hill that is before Hebron.

noctis medium et inde consurgens adprehendit ambas portae fores cum postibus suis et sera inpositasque umeris porta-

16:4 And it came to pass afterward, that he loved a woman uit ad uerticem montis qui respicit Hebron 16:4 post haec

in the valley of Sorek, whose name was Delilah.

amauit mulierem quae habitabat in ualle Sorech et uocaba-

16:5 And the lords of the Philistines came up unto her, tur Dalila 16:5 ueneruntque ad eam principes Philisthino-

and said unto her, Entice him, and see wherein his great strength lieth, and by what means we may prevail against him, that we may bind him to afflict him; and we will give thee every one of us eleven hundred pieces of silver.

rum atque dixerunt decipe eum et disce ab illo in quo tantam habeat fortitudinem et quomodo eum superare ualeamus et uinctum adfligere quod si feceris dabimus tibi singuli

16:6 And Delilah said to Samson, Tell me, I pray thee, mille centum argenteos 16:6 locuta est ergo Dalila ad Sam-

wherein thy great strength lieth, and wherewith thou mightest be bound to afflict thee.

son dic mihi obsecro in quo sit tua maxima fortitudo et quid

16:7 And Samson said unto her, If they bind me with seven sit quo ligatus erumpere nequeas 16:7 cui respondit Sam-

green withs that were never dried, then shall I be weak, and be as another man.

son si septem neruicis funibus necdum siccis et adhuc hu-

16:8 Then the lords of the Philistines brought up to her mentibus ligatus fuero infirmus ero ut ceteri homines 16:8

seven green withs which had not been dried, and she bound him with them.

adtuleruntque ad eam satrapae Philisthinorum septem funes

16:9 Now there were men lying in wait, abiding with her ut dixerat quibus uinxit eum 16:9 latentibus apud se insi-

in the chamber. And she said unto him, The Philistines be upon thee, Samson. And he brake the withs, as a thread of tow is broken when it toucheth the fire. So his strength was not known.

diis et in cubiculo finem rei expectantibus clamauitque ad eum Philisthim super te Samson qui rupit uincula quomodo si rumpat quis filum de stuppae tortum putamine cum odorem ignis acceperit et non est cognitum in quo esset fortitudo

16:10 And Delilah said unto Samson, Behold, thou hast eius 16:10 dixitque ad eum Dalila ecce inlusisti mihi et fal-

mocked me, and told me lies: now tell me, I pray thee, wherewith thou mightest be bound.

sum locutus es saltim nunc indica quo ligari debeas 16:11

16:11 And he said unto her, If they bind me fast with new ropes that never were occupied, then shall I be weak, and be as another man.

cui ille respondit si ligatus fuero nouis funibus qui numquam fuerunt in opere infirmus ero et aliorum hominum similis

16:12 Delilah therefore took new ropes, and bound him 16:12 quibus rursum Dalila uinxit eum et clamauit Philist-

therewith, and said unto him, The Philistines be upon thee, Samson. And there were liers in wait abiding in the cham-

ber. And he brake them from off his arms like a thread. him super te Samson in cubiculo insidiis praeparatis qui ita 16:13 And Delilah said unto Samson, Hitherto thou hast rupit uincula quasi fila telarum 16:13 dixitque Dalila rur-

mocked me, and told me lies: tell me wherewith thou mightest be bound. And he said unto her, If thou weavest

the seven locks of my head with the web. sum ad eum usquequo decipis me et falsum loqueris ostende

Liber Iudicum

quo uinciri debeas si inquit septem crines capitis mei cum licio plexueris et clauum his circumligatum terrae fixeris infirmus ero 16:14 quod cum fecisset Dalila dixit ad eum Phi- 16:14 And she fastened it with the pin, and said unto him,

The Philistines be upon thee, Samson. And he awaked out of his sleep, and went away with the pin of the beam, and

listhim super te Samson qui consurgens de somno extraxit with the web. clauum cum crinibus et licio 16:15 dixitque ad eum Dalila 16:15 And she said unto him, How canst thou say, I love

thee, when thine heart is not with me? thou hast mocked me these three times, and hast not told me wherein thy

quomodo dicis quod ames me cum animus tuus non sit me- great strength lieth. cum per tres uices mentitus es mihi et noluisti dicere in quo sit tua maxima fortitudo 16:16 cumque molesta ei esset et 16:16 And it came to pass, when she pressed him daily

with her words, and urged him, so that his soul was vexed unto death;

per multos dies iugiter adhereret spatium ad quietem non tribuens defecit anima eius et ad mortem usque lassata est 16:17 tunc aperiens ueritatem rei dixit ad eam ferrum num- 16:17 That he told her all his heart, and said unto her,

There hath not come a razor upon mine head; for I have been a Nazarite unto God from my mother’s womb: if I be shaven, then my strength will go from me, and I shall become weak, and be like any other man.

quam ascendit super caput meum quia nazareus id est consecratus Deo sum de utero matris meae si rasum fuerit caput meum recedet a me fortitudo mea et deficiam eroque ut ceteri homines 16:18 uidens illa quod confessus ei esset om- 16:18 And when Delilah saw that he had told her all his

heart, she sent and called for the lords of the Philistines, saying, Come up this once, for he hath shewed me all his heart. Then the lords of the Philistines came up unto her, and brought money in their hand.

nem animum suum misit ad principes Philisthinorum atque mandauit ascendite adhuc semel quia nunc mihi aperuit cor suum qui ascenderunt adsumpta pecunia quam promiserant 16:19 at illa dormire eum fecit super genua sua et in sinu 16:19 And she made him sleep upon her knees; and she

called for a man, and she caused him to shave off the seven locks of his head; and she began to afflict him, and his

suo reclinare caput uocauitque tonsorem et rasit septem cri- strength went from him. nes eius et coepit abicere eum et a se repellere statim enim ab eo fortitudo discessit 16:20 dixitque Philisthim super te 16:20 And she said, The Philistines be upon thee, Samson.

And he awoke out of his sleep, and said, I will go out as at other times before, and shake myself. And he wist not

Samson qui de somno consurgens dixit in animo suo egre- that the LORD was departed from him. diar sicut ante feci et me excutiam nesciens quod Dominus recessisset ab eo 16:21 quem cum adprehendissent Philist- 16:21 But the Philistines took him, and put out his eyes,

and brought him down to Gaza, and bound him with fetters of brass; and he did grind in the prison house.

him statim eruerunt oculos eius et duxerunt Gazam uinctum catenis et clausum in carcere molere fecerunt 16:22 iam- 16:22 Howbeit the hair of his head began to grow again

after he was shaven.

que capilli eius renasci coeperant 16:23 et principes Philis- 16:23 Then the lords of the Philistines gathered them to-

gether for to offer a great sacrifice unto Dagon their god, and to rejoice: for they said, Our god hath delivered Sam-

thinorum conuenerunt in unum ut immolarent hostias mag- son our enemy into our hand. nificas Dagon deo suo et epularentur dicentes tradidit deus noster inimicum nostrum Samson in manus nostras 16:24 16:24 And when the people saw him, they praised their

god: for they said, Our god hath delivered into our hands our enemy, and the destroyer of our country, which slew

quod etiam populus uidens laudabat deum suum eademque many of us. dicebat tradidit deus noster in manus nostras aduersarium

Biblia Sacra Vulgata

16:25 And it came to pass, when their hearts were merry, qui deleuit terram nostram et occidit plurimos 16:25 lae-

that they said, Call for Samson, that he may make us sport. And they called for Samson out of the prison house; and

he made them sport: and they set him between the pillars. tantesque per conuiuia sumptis iam epulis praeceperunt ut uocaretur Samson et ante eos luderet qui adductus de carcere ludebat ante eos feceruntque eum stare inter duas co-

16:26 And Samson said unto the lad that held him by the lumnas 16:26 qui dixit puero regenti gressus suos dimitte

hand, Suffer me that I may feel the pillars whereupon the house standeth, that I may lean upon them.

me ut tangam columnas quibus omnis inminet domus ut rec-

16:27 Now the house was full of men and women; and all liner super eas et paululum requiescam 16:27 domus autem

the lords of the Philistines were there; and there were upon the roof about three thousand men and women, that beheld

while Samson made sport. plena erat uirorum ac mulierum et erant ibi omnes princi-

pes Philisthinorum ac de tecto et solario circiter tria milia

16:28 And Samson called unto the LORD, and said, O utriusque sexus spectabant ludentem Samson 16:28 at ille

Lord God, remember me, I pray thee, and strengthen me, I pray thee, only this once, O God, that I may be at once

avenged of the Philistines for my two eyes. inuocato Domino ait Domine Deus memento mei et redde nunc mihi pristinam fortitudinem Deus meus ut ulciscar me de hostibus meis et pro amissione duorum luminum unam

16:29 And Samson took hold of the two middle pillars ultionem recipiam 16:29 et adprehendens ambas columnas

upon which the house stood, and on which it was borne up, of the one with his right hand, and of the other with

his left. quibus innitebatur domus alteramque earum dextera et al-

16:30 And Samson said, Let me die with the Philistines. teram leua tenens 16:30 ait moriatur anima mea cum Phi-And he bowed himself with all his might; and the house fell upon the lords, and upon all the people that were therein. So the dead which he slew at his death were more than they which he slew in his life.

listhim concussisque fortiter columnis cecidit domus super omnes principes et ceteram multitudinem quae ibi erat multoque plures interfecit moriens quam ante uiuus occiderat

16:31 Then his brethren and all the house of his father 16:31 descendentes autem fratres eius et uniuersa cognatio

came down, and took him, and brought him up, and buried him between Zorah and Eshtaol in the buryingplace of

Manoah his father. And he judged Israel twenty years. tulerunt corpus eius et sepelierunt inter Saraa et Esthaol in sepulchro patris Manue iudicauitque Israhel uiginti annis

17:1 And there was a man of mount Ephraim, whose name 17:1 fuit eo tempore uir quidam de monte Ephraim nomine

was Micah.

17:2 And he said unto his mother, The eleven hundred Michas 17:2 qui dixit matri suae mille centum argenteos

shekels of silver that were taken from thee, about which thou cursedst, and spakest of also in mine ears, behold, the silver is with me; I took it. And his mother said, Blessed be thou of the LORD, my son.

quos separaueras tibi et super quibus me audiente iuraueras ecce ego habeo et apud me sunt cui illa respondit benedictus

17:3 And when he had restored the eleven hundred shekels filius meus Domino 17:3 reddidit ergo eos matri suae quae of silver to his mother, his mother said, I had wholly dedicated the silver unto the LORD from my hand for my son, to make a graven image and a molten image: now therefore I will restore it unto thee.

dixerat ei consecraui et uoui argentum hoc Domino ut de manu mea suscipiat filius meus et faciat sculptile atque con-

17:4 Yet he restored the money unto his mother; and his flatile et nunc trado illud tibi 17:4 reddidit igitur matri suae

mother took two hundred shekels of silver, and gave them to the founder, who made thereof a graven image and a

molten image: and they were in the house of Micah. quae tulit ducentos argenteos et dedit eos argentario ut face-

ret ex eis sculptile atque conflatile quod fuit in domo Micha

17:5 And the man Micah had an house of gods, and made 17:5 qui aediculam quoque in ea Deo separauit et fecit ephod

an ephod, and teraphim, and consecrated one of his sons, who became his priest.

Liber Iudicum

ac therafin id est uestem sacerdotalem et idola impleuitque unius filiorum suorum manum et factus est ei sacerdos 17:6 17:6 In those days there was no king in Israel, but every

man did that which was right in his own eyes.

in diebus illis non erat rex in Israhel sed unusquisque quod sibi rectum uidebatur hoc faciebat 17:7 fuit quoque alter 17:7 And there was a young man out of Bethlehemjudah

of the family of Judah, who was a Levite, and he sojourned there.

adulescens de Bethleem Iuda et cognatione eius eratque ipse Leuites et habitabat ibi 17:8 egressusque de ciuitate Beth- 17:8 And the man departed out of the city from Bethle-

hemjudah to sojourn where he could find a place: and he came to mount Ephraim to the house of Micah, as he jour-

leem peregrinari uoluit ubicumque sibi commodum reppe- neyed. risset cumque uenisset in monte Ephraim iter faciens et declinasset parumper in domum Micha 17:9 interrogatus est 17:9 And Micah said unto him, Whence comest thou? And

he said unto him, I am a Levite of Bethlehemjudah, and I go to sojourn where I may find a place.

ab eo unde uenis qui respondit Leuita sum de Bethleem Iuda et uado ut habitem ubi potuero et utile mihi esse perspexero 17:10 mane inquit apud me et esto mihi parens ac sacerdos 17:10 And Micah said unto him, Dwell with me, and be

unto me a father and a priest, and I will give thee ten shekels of silver by the year, and a suit of apparel, and

daboque tibi per annos singulos decem argenteos ac uestem thy victuals. So the Levite went in. duplicem et quae ad uictum necessaria sunt 17:11 adquie- 17:11 And the Levite was content to dwell with the man;

and the young man was unto him as one of his sons.

uit et mansit apud hominem fuitque illi quasi unus de filiis 17:12 impleuitque Micha manum eius et habuit apud se pue- 17:12 And Micah consecrated the Levite; and the young

man became his priest, and was in the house of Micah.

rum sacerdotem 17:13 nunc scio dicens quod bene mihi fa- 17:13 Then said Micah, Now know I that the LORD will

do me good, seeing I have a Levite to my priest.

ciat Deus habenti leuitici generis sacerdotem

18:1 in diebus illis non erat rex in Israhel et tribus Dan 18:1 In those days there was no king in Israel: and in those days the tribe of the Danites sought them an inheritance to dwell in; for unto that day all their inheritance had not

quaerebat possessionem sibi ut habitaret in ea usque ad illum fallen unto them among the tribes of Israel. enim diem inter ceteras tribus sortem non acceperat 18:2 18:2 And the children of Dan sent of their family five men

from their coasts, men of valour, from Zorah, and from Eshtaol, to spy out the land, and to search it; and they said unto them, Go, search the land: who when they came to mount Ephraim, to the house of Micah, they lodged there.

miserunt igitur filii Dan stirpis et familiae suae quinque uiros fortissimos de Saraa et Esthaol ut explorarent terram et diligenter inspicerent dixeruntque eis ite et considerate terram qui cum pergentes uenissent in montem Ephraim et intrassent domum Micha requieuerunt ibi 18:3 et agnoscen- 18:3 When they were by the house of Micah, they knew

the voice of the young man the Levite: and they turned in thither, and said unto him, Who brought thee hither? and

tes uocem adulescentis Leuitae utentesque illius diuersorio what makest thou in this place? and what hast thou here? dixerunt ad eum quis te huc adduxit quid hic agis quam ob causam huc uenire uoluisti 18:4 qui respondit eis haec et 18:4 And he said unto them, Thus and thus dealeth Micah

with me, and hath hired me, and I am his priest.

haec praestitit mihi Michas et me mercede conduxit ut sim ei sacerdos 18:5 rogaueruntque eum ut consuleret Dominum 18:5 And they said unto him, Ask counsel, we pray thee,

of God, that we may know whether our way which we go shall be prosperous.

et scire possent an prospero itinere pergerent et res haberet effectum 18:6 qui respondit eis ite cum pace Dominus re- 18:6 And the priest said unto them, Go in peace: before

the LORD is your way wherein ye go.

Biblia Sacra Vulgata

18:7 Then the five men departed, and came to Laish, and spicit uiam uestram et iter quo pergitis 18:7 euntes itaque

saw the people that were therein, how they dwelt careless, after the manner of the Zidonians, quiet and secure; and there was no magistrate in the land, that might put them to shame in any thing; and they were far from the Zidonians, and had no business with any man.

quinque uiri uenerunt Lais uideruntque populum habitantem in ea absque ullo timore iuxta Sidoniorum consuetudinem securum et quietum nullo eis penitus resistente magnarumque opum et procul a Sidone atque a cunctis hominibus se-

18:8 And they came unto their brethren to Zorah and Esh- paratum 18:8 reuersique ad fratres suos in Saraa et Esthaol

taol: and their brethren said unto them, What say ye?

18:9 And they said, Arise, that we may go up against them: et quid egissent sciscitantibus responderunt 18:9 surgite et

for we have seen the land, and, behold, it is very good: and are ye still? be not slothful to go, and to enter to possess

the land. ascendamus ad eos uidimus enim terram ualde opulentam

et uberem nolite neglegere nolite cessare eamus et posside-

18:10 When ye go, ye shall come unto a people secure, amus eam nullus erit labor 18:10 intrabimus ad securos in

and to a large land: for God hath given it into your hands; a place where there is no want of any thing that is in the

earth. regionem latissimam tradetque nobis Dominus locum in quo 18:11 And there went from thence of the family of the nullius rei est penuria eorum quae gignuntur in terra 18:11

Danites, out of Zorah and out of Eshtaol, six hundred men appointed with weapons of war.

profecti igitur sunt de cognatione Dan id est de Saraa et Est-

18:12 And they went up, and pitched in Kirjathjearim, in haol sescenti uiri accincti armis bellicis 18:12 ascendentes-

Judah: wherefore they called that place Mahanehdan unto this day: behold, it is behind Kirjathjearim.

que manserunt in Cariathiarim Iudae qui locus ex eo tempore castrorum Dan nomen accepit et est post tergum Ca-

18:13 And they passed thence unto mount Ephraim, and riathiarim 18:13 inde transierunt in montem Ephraim cum-

came unto the house of Micah.

18:14 Then answered the five men that went to spy out que uenissent ad domum Micha 18:14 dixerunt quinque uiri

the country of Laish, and said unto their brethren, Do ye know that there is in these houses an ephod, and teraphim, and a graven image, and a molten image? now therefore consider what ye have to do.

qui prius missi fuerant ad considerandam terram Lais ceteris fratribus suis nostis quod in domibus istis sit ephod et therafin et sculptile atque conflatile uidete quid uobis placeat

18:15 And they turned thitherward, and came to the house 18:15 et cum paululum declinassent ingressi sunt domum ad-

of the young man the Levite, even unto the house of Micah, and saluted him.

ulescentis Leuitae qui erat in domo Micha salutaueruntque

18:16 And the six hundred men appointed with their eum uerbis pacificis 18:16 sescenti autem uiri ita ut erant

weapons of war, which were of the children of Dan, stood by the entering of the gate.

armati stabant ante ostium 18:17 at illi qui ingressi fuerant

18:17 And the five men that went to spy out the land went up, and came in thither, and took the graven image, and the ephod, and the teraphim, and the molten image: and the priest stood in the entering of the gate with the six hundred men that were appointed with weapons of war.

domum iuuenis sculptile et ephod et therafin atque conflatile tollere nitebantur et sacerdos stabat ante ostium sescen-

18:18 And these went into Micah’s house, and fetched the tis uiris fortissimis haut procul expectantibus 18:18 tulerunt

carved image, the ephod, and the teraphim, and the molten image. Then said the priest unto them, What do ye?

igitur qui intrauerant sculptile ephod et idola atque confla-

18:19 And they said unto him, Hold thy peace, lay thine tile quibus dixit sacerdos quid facitis 18:19 cui responde-

hand upon thy mouth, and go with us, and be to us a father and a priest: is it better for thee to be a priest unto the house of one man, or that thou be a priest unto a tribe and a family in Israel?

runt tace et pone digitum super os tuum uenique nobiscum ut habeamus te patrem et sacerdotem quid tibi melius est ut sis sacerdos in domo unius uiri an in una tribu et familia in

Liber Iudicum

Israhel 18:20 quod cum audisset adquieuit sermonibus eo- 18:20 And the priest’s heart was glad, and he took the

ephod, and the teraphim, and the graven image, and went in the midst of the people.

rum et tulit ephod et idola ac sculptile et cum eis profectus est 18:21 qui cum pergerent et ante se ire fecissent paruu- 18:21 So they turned and departed, and put the little ones

and the cattle and the carriage before them.

los et iumenta et omne quod erat pretiosum 18:22 iamque 18:22 And when they were a good way from the house of

Micah, the men that were in the houses near to Micah’s house were gathered together, and overtook the children

a domo Michae essent procul uiri qui habitabant in aedibus of Dan. Michae conclamantes secuti sunt 18:23 et post tergum cla- 18:23 And they cried unto the children of Dan. And they

turned their faces, and said unto Micah, What aileth thee, that thou comest with such a company?

mare coeperunt qui cum respexissent dixerunt ad Micham quid tibi uis cur clamas 18:24 qui respondit deos meos quos 18:24 And he said, Ye have taken away my gods which I

made, and the priest, and ye are gone away: and what have I more? and what is this that ye say unto me, What aileth

mihi feci tulistis et sacerdotem et omnia quae habeo et dici- thee? tis quid tibi est 18:25 dixeruntque ei filii Dan caue ne ultra 18:25 And the children of Dan said unto him, Let not thy

voice be heard among us, lest angry fellows run upon thee, and thou lose thy life, with the lives of thy household.

loquaris ad nos et ueniant ad te uiri animo concitati et ipse cum omni domo tua pereas 18:26 et sic coepto itinere per- 18:26 And the children of Dan went their way: and when

Micah saw that they were too strong for him, he turned and went back unto his house.

rexerunt uidens autem Micha quod fortiores se essent reuersus est in domum suam 18:27 sescenti autem uiri tulerunt 18:27 And they took the things which Micah had made,

and the priest which he had, and came unto Laish, unto a people that were at quiet and secure: and they smote them

sacerdotem et quae supra diximus ueneruntque in Lais ad with the edge of the sword, and burnt the city with fire. populum quiescentem atque securum et percusserunt eos in ore gladii urbemque incendio tradiderunt 18:28 nullo peni- 18:28 And there was no deliverer, because it was far from

Zidon, and they had no business with any man; and it was in the valley that lieth by Bethrehob. And they built a city,

tus ferente praesidium eo quod procul habitarent a Sidone and dwelt therein. et cum nullo hominum haberent quicquam societatis ac negotii erat autem ciuitas sita in regione Roob quam rursum extruentes habitauerunt in ea 18:29 uocato nomine ciuitatis 18:29 And they called the name of the city Dan, after the

name of Dan their father, who was born unto Israel: howbeit the name of the city was Laish at the first.

Dan iuxta uocabulum patris sui quem genuerat Israhel quae prius Lais dicebatur 18:30 posueruntque sibi sculptile et Io- 18:30 And the children of Dan set up the graven image:

and Jonathan, the son of Gershom, the son of Manasseh, he and his sons were priests to the tribe of Dan until the

nathan filium Gersan filii Mosi ac filios eius sacerdotes in day of the captivity of the land. tribu Dan usque ad diem captiuitatis suae 18:31 mansitque 18:31 And they set them up Micah’s graven image, which

he made, all the time that the house of God was in Shiloh.

apud eos idolum Michae omni tempore quo fuit domus Dei in Silo in diebus illis non erat rex in Israhel

19:1 fuit quidam uir Leuites habitans in latere montis Eph- 19:1 And it came to pass in those days, when there was no

king in Israel, that there was a certain Levite sojourning on the side of mount Ephraim, who took to him a concubine

raim qui accepit uxorem de Bethleem Iuda 19:2 quae reli- out of Bethlehemjudah.

19:2 And his concubine played the whore against him, and went away from him unto her father’s house to Bethlehemjudah, and was there four whole months.

quit eum et reuersa est in domum patris sui Bethleem mansitque apud eum quattuor mensibus 19:3 secutusque est eam 19:3 And her husband arose, and went after her, to speak

friendly unto her, and to bring her again, having his servant with him, and a couple of asses: and she brought him into her father’s house: and when the father of the damsel saw him, he rejoiced to meet him.

uir suus uolens ei reconciliari atque blandiri et secum reducere habens in comitatu puerum et duos asinos quae su-

Biblia Sacra Vulgata

scepit eum et introduxit in domum patris sui quod cum au-

19:4 And his father in law, the damsel’s father, retained disset socer eius eumque uidisset occurrit ei laetus 19:4 et

him; and he abode with him three days: so they did eat and drink, and lodged there.

amplexatus est hominem mansitque gener in domo soceri

19:5 And it came to pass on the fourth day, when they tribus diebus comedens cum eo et bibens familiariter 19:5

arose early in the morning, that he rose up to depart: and the damsel’s father said unto his son in law, Comfort thine

heart with a morsel of bread, and afterward go your way. die autem quarto de nocte consurgens proficisci uoluit quem

tenuit socer et ait ad eum gusta prius pauxillum panis et con-

19:6 And they sat down, and did eat and drink both of forta stomachum et sic proficisceris 19:6 sederuntque simul

them together: for the damsel’s father had said unto the man, Be content, I pray thee, and tarry all night, and let

thine heart be merry. et comederunt ac biberunt dixitque pater puellae ad gene-

rum suum quaeso te ut hodie hic maneas pariterque laetemur

19:7 And when the man rose up to depart, his father in law 19:7 at ille consurgens coepit uelle proficisci et nihilominus

urged him: therefore he lodged there again.

19:8 And he arose early in the morning on the fifth day to obnixe eum socer tenuit et apud se fecit manere 19:8 mane

depart; and the damsel’s father said, Comfort thine heart, I pray thee. And they tarried until afternoon, and they did

eat both of them. facto parabat Leuites iter cui rursum socer oro te inquit ut

paululum cibi capias et adsumptis uiribus donec increscat

19:9 And when the man rose up to depart, he, and his con- dies postea proficiscaris comederunt ergo simul 19:9 sur-

cubine, and his servant, his father in law, the damsel’s father, said unto him, Behold, now the day draweth toward evening, I pray you tarry all night: behold, the day groweth to an end, lodge here, that thine heart may be merry; and to morrow get you early on your way, that thou mayest go home.

rexitque adulescens ut pergeret cum uxore sua et puero cui rursum locutus est socer considera quod dies ad occasum decliuior sit et propinquet ad uesperum mane apud me etiam hodie et duc laetum diem et cras proficisceris ut uadas in do-

19:10 But the man would not tarry that night, but he rose mum tuam 19:10 noluit gener adquiescere sermonibus eius

up and departed, and came over against Jebus, which is Jerusalem; and there were with him two asses saddled, his

concubine also was with him. sed statim perrexit et uenit contra Iebus quae altero nomine

uocabatur Hierusalem ducens secum duos asinos onustos et

19:11 And when they were by Jebus, the day was far spent; concubinam 19:11 iamque aderant iuxta Iebus et dies mu-

and the servant said unto his master, Come, I pray thee, and let us turn in into this city of the Jebusites, and lodge

in it. tabatur in noctem dixitque puer ad dominum suum ueni ob-

secro declinemus ad urbem Iebuseorum et maneamus in ea

19:12 And his master said unto him, We will not turn aside 19:12 cui respondit dominus non ingrediar oppidum gentis

hither into the city of a stranger, that is not of the children of Israel; we will pass over to Gibeah.

alienae quae non est de filiis Israhel sed transibo usque Ga-

19:13 And he said unto his servant, Come, and let us draw baa 19:13 et cum illuc peruenero manebimus in ea aut certe

near to one of these places to lodge all night, in Gibeah, or in Ramah.

in urbe Rama 19:14 transierunt igitur Iebus et coeptum car-

19:14 And they passed on and went their way; and the sun went down upon them when they were by Gibeah, which belongeth to Benjamin.

pebant iter occubuitque eis sol iuxta Gabaa quae est in tribu

19:15 And they turned aside thither, to go in and to lodge Beniamin 19:15 deuerteruntque ad eam ut manerent ibi quo

in Gibeah: and when he went in, he sat him down in a street of the city: for there was no man that took them into

his house to lodging. cum intrassent sedebant in platea ciuitatis et nullus eos re-19:16 And, behold, there came an old man from his work cipere uolebat hospitio 19:16 et ecce apparuit homo senex

out of the field at even, which was also of mount Ephraim; and he sojourned in Gibeah: but the men of the place were

Benjamites. reuertens de agro et de opere suo uespere qui et ipse erat

Liber Iudicum

de monte Ephraim et peregrinus habitabat in Gabaa homines autem regionis illius erant filii Iemini 19:17 eleuatisque 19:17 And when he had lifted up his eyes, he saw a way-

faring man in the street of the city: and the old man said, Whither goest thou? and whence comest thou?

oculis uidit senex sedentem hominem cum sarcinulis suis in platea ciuitatis et dixit ad eum unde uenis et quo uadis 19:18 19:18 And he said unto him, We are passing from Bethle-

hemjudah toward the side of mount Ephraim; from thence am I: and I went to Bethlehemjudah, but I am now going to the house of the LORD; and there is no man that receiveth me to house.

qui respondit ei profecti sumus de Bethleem Iuda et pergimus ad locum nostrum qui est in latere montis Ephraim unde ieramus Bethleem et nunc uadimus ad domum Dei nullusque nos sub tectum suum uult recipere 19:19 habentes paleas et 19:19 Yet there is both straw and provender for our asses;

and there is bread and wine also for me, and for thy handmaid, and for the young man which is with thy servants:

faenum in asinorum pabulum et panem ac uinum in meos there is no want of any thing. et ancillae tuae usus et pueri qui mecum est nulla re indigemus nisi hospitio 19:20 cui respondit senex pax tecum sit 19:20 And the old man said, Peace be with thee; howso-

ever let all thy wants lie upon me; only lodge not in the street.

ego praebebo omnia quae necessaria sunt tantum quaeso ne in platea maneas 19:21 introduxitque eum in domum suam 19:21 So he brought him into his house, and gave proven-

der unto the asses: and they washed their feet, and did eat and drink.

et pabulum asinis praebuit ac postquam lauerunt pedes suos recepit eos in conuiuium 19:22 illis epulantibus et post la- 19:22 Now as they were making their hearts merry, be-

hold, the men of the city, certain sons of Belial, beset the house round about, and beat at the door, and spake to the master of the house, the old man, saying, Bring forth the man that came into thine house, that we may know him.

borem itineris cibo ac potu reficientibus corpora uenerunt uiri ciuitatis illius filii Belial id est absque iugo et circumdantes domum senis fores pulsare coeperunt clamantes ad dominum domus atque dicentes educ uirum qui ingressus

est domum tuam ut abutamur eo 19:23 egressusque est ad 19:23 And the man, the master of the house, went out unto them, and said unto them, Nay, my brethren, nay, I pray you, do not so wickedly; seeing that this man is come into

eos senex et ait nolite fratres nolite facere malum hoc quia mine house, do not this folly. ingressus est homo hospitium meum et cessate ab hac stultitia 19:24 habeo filiam uirginem et hic homo habet concu- 19:24 Behold, here is my daughter a maiden, and his con-

cubine; them I will bring out now, and humble ye them, and do with them what seemeth good unto you: but unto

binam educam eas ad uos ut humilietis eas et uestram libi- this man do not so vile a thing. dinem conpleatis tantum obsecro ne scelus hoc contra naturam operemini in uirum 19:25 nolebant adquiescere ser- 19:25 But the men would not hearken to him: so the man

took his concubine, and brought her forth unto them; and they knew her, and abused her all the night until the morn-

monibus eius quod cernens homo eduxit ad eos concubinam ing: and when the day began to spring, they let her go. suam et eis tradidit inludendam qua cum tota nocte abusi essent dimiserunt eam mane 19:26 at mulier recedentibus 19:26 Then came the woman in the dawning of the day,

and fell down at the door of the man’s house where her lord was, till it was light.

tenebris uenit ad ostium domus ubi manebat dominus suus et ibi corruit 19:27 mane facto surrexit homo et aperuit os- 19:27 And her lord rose up in the morning, and opened

the doors of the house, and went out to go his way: and, behold, the woman his concubine was fallen down at the

tium ut coeptam expleret uiam et ecce concubina eius iace- door of the house, and her hands were upon the threshold. bat ante ostium sparsis in limine manibus 19:28 cui ille pu- 19:28 And he said unto her, Up, and let us be going. But

none answered. Then the man took her up upon an ass, and the man rose up, and gat him unto his place.

Biblia Sacra Vulgata

tans eam quiescere loquebatur surge ut ambulemus qua nihil respondente intellegens quod erat tulit eam et inposuit asino

19:29 And when he was come into his house, he took a reuersusque est in domum suam 19:29 quam cum esset in-

knife, and laid hold on his concubine, and divided her, together with her bones, into twelve pieces, and sent her

into all the coasts of Israel. gressus arripuit gladium et cadauer uxoris cum ossibus suis

in duodecim partes ac frusta concidens misit in omnes termi-

19:30 And it was so, that all that saw it said, There was nos Israhel 19:30 quod cum uidissent singuli conclamabant

no such deed done nor seen from the day that the children of Israel came up out of the land of Egypt unto this day:

consider of it, take advice, and speak your minds. numquam res talis facta est in Israhel ex eo die quo ascenderunt patres nostri de Aegypto usque in praesens tempus ferte sententiam et in commune decernite quid facto opus sit

20:1 Then all the children of Israel went out, and the con- 20:1 egressi sunt itaque omnes filii Israhel et pariter con-

gregation was gathered together as one man, from Dan even to Beersheba, with the land of Gilead, unto the

LORD in Mizpeh. gregati quasi uir unus de Dan usque Bersabee et terra Galaad

20:2 And the chief of all the people, even of all the tribes ad Dominum in Maspha 20:2 omnesque anguli populorum of Israel, presented themselves in the assembly of the people of God, four hundred thousand footmen that drew

sword. et cunctae tribus Israhel in ecclesiam populi Dei conuene-20:3 (Now the children of Benjamin heard that the chil- runt quadringenta milia peditum pugnatorum 20:3 nec la-

dren of Israel were gone up to Mizpeh.) Then said the children of Israel, Tell us, how was this wickedness?

tuit filios Beniamin quod ascendissent filii Israhel in Maspha interrogatusque Leuita maritus mulieris interfectae quo-

20:4 And the Levite, the husband of the woman that was modo tantum scelus perpetratum esset 20:4 respondit ueni

slain, answered and said, I came into Gibeah that belongeth to Benjamin, I and my concubine, to lodge.

in Gabaa Beniamin cum uxore mea illucque deuerti 20:5 et

20:5 And the men of Gibeah rose against me, and beset the house round about upon me by night, and thought to have slain me: and my concubine have they forced, that she is dead.

ecce homines ciuitatis illius circumdederunt nocte domum in qua manebam uolentes me occidere et uxorem meam in-

20:6 And I took my concubine, and cut her in pieces, and credibili libidinis furore uexantes denique mortua est 20:6

sent her throughout all the country of the inheritance of Israel: for they have committed lewdness and folly in Israel.

quam arreptam in frusta concidi misique partes in omnes terminos possessionis uestrae quia numquam tantum nefas et

20:7 Behold, ye are all children of Israel; give here your tam grande piaculum factum est in Israhel 20:7 adestis om-

advice and counsel.

20:8 And all the people arose as one man, saying, We will nes filii Israhel decernite quid facere debeatis 20:8 stans-

not any of us go to his tent, neither will we any of us turn into his house.

que omnis populus quasi unius hominis sermone respondit non recedemus in tabernacula nostra nec suam quisquam in-

20:9 But now this shall be the thing which we will do to trabit domum 20:9 sed hoc contra Gabaa in commune fa-Gibeah; we will go up by lot against it; 20:10 And we will take ten men of an hundred throughout ciemus 20:10 decem uiri eligantur e centum ex omnibus all the tribes of Israel, and an hundred of a thousand, and a thousand out of ten thousand, to fetch victual for the people, that they may do, when they come to Gibeah of Benjamin, according to all the folly that they have wrought in Israel.

tribubus Israhel et centum de mille et mille de decem milibus ut conportent exercitui cibaria et possimus pugnantes contra Gabaa Beniamin reddere ei pro scelere quod mere-

20:11 So all the men of Israel were gathered against the tur 20:11 conuenitque uniuersus Israhel ad ciuitatem quasi

city, knit together as one man.

Liber Iudicum

unus homo eadem mente unoque consilio 20:12 et miserunt 20:12 And the tribes of Israel sent men through all the tribe of Benjamin, saying, What wickedness is this that is done among you?

nuntios ad omnem tribum Beniamin qui dicerent cur tantum nefas in uobis reppertum est 20:13 tradite homines de Ga- 20:13 Now therefore deliver us the men, the children of

Belial, which are in Gibeah, that we may put them to death, and put away evil from Israel. But the children of Benjamin would not hearken to the voice of their brethren the children of Israel.

baa qui hoc flagitium perpetrarunt ut moriantur et auferatur malum de Israhel qui noluerunt fratrum suorum filiorum Israhel audire mandatum 20:14 sed ex cunctis urbibus quae 20:14 But the children of Benjamin gathered themselves

together out of the cities unto Gibeah, to go out to battle against the children of Israel.

suae sortis erant conuenerunt in Gabaa ut illis ferrent auxilium et contra uniuersum Israhel populum dimicarent 20:15 20:15 And the children of Benjamin were numbered at that

time out of the cities twenty and six thousand men that drew sword, beside the inhabitants of Gibeah, which were

inuentique sunt uiginti quinque milia de Beniamin educen- numbered seven hundred chosen men. tium gladium praeter habitatores Gabaa 20:16 qui septin- 20:16 Among all this people there were seven hundred

chosen men lefthanded; every one could sling stones at an hair breadth, and not miss.

genti erant uiri fortissimi ita sinistra ut dextra proeliantes et sic fundis ad certum iacientes lapides ut capillum quoque possent percutere et nequaquam in alteram partem ictus lapidis deferretur 20:17 uirorum quoque Israhel absque filiis 20:17 And the men of Israel, beside Benjamin, were num-

bered four hundred thousand men that drew sword: all these were men of war.

Beniamin inuenta sunt quadringenta milia educentium gladios et paratorum ad pugnam 20:18 qui surgentes uenerunt 20:18 And the children of Israel arose, and went up to the

house of God, and asked counsel of God, and said, Which of us shall go up first to the battle against the children of

in domum Dei hoc est in Silo consulueruntque eum atque Benjamin? And the LORD said, Judah shall go up first. dixerunt quis erit in exercitu nostro princeps certaminis contra filios Beniamin quibus respondit Dominus Iudas sit dux uester 20:19 statimque filii Israhel surgentes mane castra- 20:19 And the children of Israel rose up in the morning,

and encamped against Gibeah.

metati sunt iuxta Gabaa 20:20 et inde procedentes ad pu- 20:20 And the men of Israel went out to battle against Benjamin; and the men of Israel put themselves in array to fight against them at Gibeah.

gnam contra Beniamin urbem obpugnare coeperunt 20:21

20:21 And the children of Benjamin came forth out of Gibeah, and destroyed down to the ground of the Israelites that day twenty and two thousand men.

egressique filii Beniamin de Gabaa occiderunt de filiis Israhel die illo uiginti duo milia uiros 20:22 rursum filii Is- 20:22 And the people the men of Israel encouraged them-

selves, and set their battle again in array in the place where they put themselves in array the first day.

rahel et fortitudine et numero confidentes in eodem loco in quo prius certauerant aciem direxerunt 20:23 ita tamen ut 20:23 (And the children of Israel went up and wept before

the LORD until even, and asked counsel of the LORD, saying, Shall I go up again to battle against the children of Benjamin my brother? And the LORD said, Go up against him.)

prius ascenderent et flerent coram Domino usque ad noctem consulerentque eum et dicerent debeo ultra procedere ad dimicandum contra filios Beniamin fratres meos an non quibus ille respondit ascendite ad eum et inite certamen 20:24 20:24 And the children of Israel came near against the

children of Benjamin the second day.

cumque filii Israhel altero die contra Beniamin ad proelium processissent 20:25 eruperunt filii Beniamin de portis Ga- 20:25 And Benjamin went forth against them out of

Gibeah the second day, and destroyed down to the ground of the children of Israel again eighteen thousand men; all

baa et occurrentes eis tanta in illos caede baccati sunt ut these drew the sword.

Biblia Sacra Vulgata

decem et octo milia uirorum educentium gladium proster-

20:26 Then all the children of Israel, and all the people, nerent 20:26 quam ob rem omnes filii Israhel uenerunt in went up, and came unto the house of God, and wept, and sat there before the LORD, and fasted that day until even, and offered burnt offerings and peace offerings before the LORD.

domum Dei et sedentes flebant coram Domino ieiunaueruntque illo die usque ad uesperam et obtulerunt ei holocausta et

20:27 And the children of Israel enquired of the LORD, pacificas uictimas 20:27 et super statu suo interrogauerunt

(for the ark of the covenant of God was there in those days,

20:28 And Phinehas, the son of Eleazar, the son of Aaron, eo tempore ibi erat arca foederis Dei 20:28 et Finees filius

stood before it in those days,) saying, Shall I yet again go out to battle against the children of Benjamin my brother, or shall I cease? And the LORD said, Go up; for to morrow I will deliver them into thine hand.

Eleazari filii Aaron praepositus domus consuluerunt igitur Dominum atque dixerunt exire ultra debemus ad pugnam contra filios Beniamin fratres nostros an quiescere quibus ait Dominus ascendite cras enim tradam eos in manus ue-

20:29 And Israel set liers in wait round about Gibeah. stras 20:29 posueruntque filii Israhel insidias per circuitum

20:30 And the children of Israel went up against the chil- urbis Gabaa 20:30 et tertia uice sicut semel et bis contra dren of Benjamin on the third day, and put themselves in array against Gibeah, as at other times.

Beniamin exercitum produxerunt 20:31 sed et filii Benia-

20:31 And the children of Benjamin went out against the people, and were drawn away from the city; and they began to smite of the people, and kill, as at other times, in the highways, of which one goeth up to the house of God, and the other to Gibeah in the field, about thirty men of Israel.

min audacter eruperunt de ciuitate et fugientes aduersarios longius persecuti sunt ita ut uulnerarent ex eis sicut primo et secundo die et caederent per duas semitas terga uertentes quarum una ferebat in Bethel altera in Gabaa atque proster-

20:32 And the children of Benjamin said, They are smitten nerent triginta circiter uiros 20:32 putauerunt enim solito

down before us, as at the first. But the children of Israel said, Let us flee, and draw them from the city unto the

highways. eos more cedere qui fugam arte simulantes iniere consilium

ut abstraherent eos de ciuitate et quasi fugientes ad supradic-

20:33 And all the men of Israel rose up out of their place, tas semitas perducerent 20:33 omnes itaque filii Israhel sur-

and put themselves in array at Baaltamar: and the liers in wait of Israel came forth out of their places, even out of

the meadows of Gibeah. gentes de sedibus suis tetenderunt aciem in loco qui uocatur

Baalthamar insidiae quoque quae circa urbem erant paulatim

20:34 And there came against Gibeah ten thousand chosen se aperire coeperunt 20:34 et ab occidentali urbis parte pro-

men out of all Israel, and the battle was sore: but they knew not that evil was near them.

cedere sed et alia decem milia uirorum de uniuerso Israhel habitatores urbis ad certamina prouocabant ingrauatumque est bellum contra filios Beniamin et non intellexerunt quod

20:35 And the LORD smote Benjamin before Israel: and ex omni parte illis instaret interitus 20:35 percussitque eos

the children of Israel destroyed of the Benjamites that day twenty and five thousand and an hundred men: all these

drew the sword. Dominus in conspectu filiorum Israhel et interfecerunt ex

eis in illo die uiginti quinque milia et centum uiros omnes

20:36 So the children of Benjamin saw that they were bellatores et educentes gladium 20:36 filii autem Beniamin

smitten: for the men of Israel gave place to the Benjamites, because they trusted unto the liers in wait which they had

set beside Gibeah. cum se inferiores esse uidissent coeperunt fugere quod cernentes filii Israhel dederunt eis ad fugiendum locum ut ad

Liber Iudicum

praeparatas insidias deuenirent quas iuxta urbem posuerant 20:37 qui cum repente de latibulis surrexissent et Beniamin 20:37 And the liers in wait hasted, and rushed upon

Gibeah; and the liers in wait drew themselves along, and smote all the city with the edge of the sword.

terga caedentibus daret ingressi sunt ciuitatem et percusserunt eam in ore gladii 20:38 signum autem dederant filii Is- 20:38 Now there was an appointed sign between the men

of Israel and the liers in wait, that they should make a great flame with smoke rise up out of the city.

rahel his quos in insidiis conlocauerant ut postquam urbem cepissent ignem accenderent et ascendente in altum fumo captam urbem demonstrarent 20:39 quod cum cernerent filii 20:39 And when the men of Israel retired in the battle,

Benjamin began to smite and kill of the men of Israel about thirty persons: for they said, Surely they are smitten down

Israhel in ipso certamine positi putauerunt enim filii Benia- before us, as in the first battle. min eos fugere et instantius sequebantur caesis de exercitu eorum triginta uiris 20:40 et uiderent quasi columnam fumi 20:40 But when the flame began to arise up out of the

city with a pillar of smoke, the Benjamites looked behind them, and, behold, the flame of the city ascended up to

de ciuitate conscendere Beniamin quoque retro aspiciens ca- heaven. ptam cerneret ciuitatem et flammas in sublime ferri 20:41 20:41 And when the men of Israel turned again, the men of

Benjamin were amazed: for they saw that evil was come upon them.

qui prius simulauerant fugam uersa facie fortius resistebant quod cum uidissent filii Beniamin in fugam uersi sunt 20:42 20:42 Therefore they turned their backs before the men of

Israel unto the way of the wilderness; but the battle overtook them; and them which came out of the cities they

et ad uiam deserti ire coeperunt illuc quoque eos aduersariis destroyed in the midst of them. persequentibus sed et hii qui urbem succenderant occurre-

runt eis 20:43 atque ita factum est ut ex utraque parte ab 20:43 Thus they inclosed the Benjamites round about, and chased them, and trode them down with ease over against Gibeah toward the sunrising.

hostibus caederentur nec erat ulla morientium requies ceciderunt atque prostrati sunt ad orientalem plagam urbis Gabaa 20:44 fuerunt autem qui in eodem loco interfecti sunt 20:44 And there fell of Benjamin eighteen thousand men;

all these were men of valour.

decem et octo milia uirorum omnes robustissimi pugnatores 20:45 quod cum uidissent qui remanserant de Beniamin 20:45 And they turned and fled toward the wilderness unto

the rock of Rimmon: and they gleaned of them in the highways five thousand men; and pursued hard after them unto

fugerunt in solitudinem et pergebant ad petram cuius uoca- Gidom, and slew two thousand men of them. bulum est Remmon in illa quoque fuga palantes et in diuersa tendentes occiderunt quinque milia uiros et cum ultra tenderent persecuti sunt eos et interfecerunt etiam alios duo milia 20:46 et sic factum est ut omnes qui ceciderant de Beniamin 20:46 So that all which fell that day of Benjamin were

twenty and five thousand men that drew the sword; all these were men of valour.

in diuersis locis essent uiginti quinque milia pugnatores ad bella promptissimi 20:47 remanserunt itaque de omni nu- 20:47 But six hundred men turned and fled to the wilder-

ness unto the rock Rimmon, and abode in the rock Rimmon four months.

mero Beniamin qui euadere potuerant et fugere in solitudinem sescenti uiri sederuntque in petra Remmon mensibus quattuor 20:48 regressi autem filii Israhel omnes reliquias 20:48 And the men of Israel turned again upon the children

of Benjamin, and smote them with the edge of the sword, as well the men of every city, as the beast, and all that came to hand: also they set on fire all the cities that they came to.

ciuitatis a uiris usque ad iumenta gladio percusserunt cunc-

Biblia Sacra Vulgata

tasque urbes et uiculos Beniamin uorax flamma consumpsit

21:1 Now the men of Israel had sworn in Mizpeh, saying, 21:1 iurauerunt quoque filii Israhel in Maspha et dixerunt

There shall not any of us give his daughter unto Benjamin to wife.

nullus nostrum dabit filiis Beniamin de filiabus suis uxorem

21:2 And the people came to the house of God, and abode 21:2 ueneruntque omnes ad domum Dei in Silo et in conthere till even before God, and lifted up their voices, and wept sore;

spectu eius sedentes usque ad uesperam leuauerunt uocem et

21:3 And said, O LORD God of Israel, why is this come to magno ululatu coeperunt flere dicentes 21:3 quare Domine

pass in Israel, that there should be to day one tribe lacking in Israel?

Deus Israhel factum est hoc malum in populo tuo ut hodie

21:4 And it came to pass on the morrow, that the people una tribus auferretur ex nobis 21:4 altera autem die diluculo

rose early, and built there an altar, and offered burnt offerings and peace offerings.

consurgentes extruxerunt altare obtuleruntque ibi holocausta

21:5 And the children of Israel said, Who is there among et pacificas uictimas et dixerunt 21:5 quis non ascendit in

all the tribes of Israel that came not up with the congregation unto the LORD? For they had made a great oath concerning him that came not up to the LORD to Mizpeh, saying, He shall surely be put to death.

exercitu Domini de uniuersis tribubus Israhel grandi enim se iuramento constrinxerant cum essent in Maspha interfici eos

21:6 And the children of Israel repented them for Ben- qui defuissent 21:6 ductique paenitentia filii Israhel super

jamin their brother, and said, There is one tribe cut off from Israel this day.

fratre suo Beniamin coeperunt dicere ablata est una tribus

21:7 How shall we do for wives for them that remain, see- de Israhel 21:7 unde uxores accipient omnes enim in com-

ing we have sworn by the LORD that we will not give them of our daughters to wives?

mune iurauimus non daturos nos his filias nostras 21:8 id-

21:8 And they said, What one is there of the tribes of Israel that came not up to Mizpeh to the LORD? And, behold, there came none to the camp from Jabeshgilead to the assembly.

circo dixerunt quis est de uniuersis tribubus Israhel qui non ascendit ad Dominum in Maspha et ecce inuenti sunt habi-

21:9 For the people were numbered, and, behold, there tatores Iabisgalaad in illo exercitu non fuisse 21:9 eo quo-

were none of the inhabitants of Jabeshgilead there.

que tempore cum essent in Silo nullus ex eis ibi reppertus

21:10 And the congregation sent thither twelve thousand est 21:10 miserunt itaque decem milia uiros robustissimos

men of the valiantest, and commanded them, saying, Go and smite the inhabitants of Jabeshgilead with the edge of

the sword, with the women and the children. et praeceperunt eis ite et percutite habitatores Iabisgalaad in

21:11 And this is the thing that ye shall do, Ye shall utterly ore gladii tam uxores quam paruulos eorum 21:11 et hoc destroy every male, and every woman that hath lain by man.

erit quod obseruare debetis omne generis masculini et mu-

21:12 And they found among the inhabitants of lieres quae cognouerunt uiros interficite 21:12 inuentaeque

Jabeshgilead four hundred young virgins, that had known no man by lying with any male: and they brought them

unto the camp to Shiloh, which is in the land of Canaan. sunt de Iabisgalaad quadringentae uirgines quae nescierunt

uiri torum et adduxerunt eas in castra in Silo in terra Cha-

21:13 And the whole congregation sent some to speak to naan 21:13 miseruntque nuntios ad filios Beniamin qui er-

the children of Benjamin that were in the rock Rimmon, and to call peaceably unto them.

ant in petra Remmon et praeceperunt eis ut eos in pace sus-

21:14 And Benjamin came again at that time; and they ciperent 21:14 ueneruntque filii Beniamin in illo tempore et

gave them wives which they had saved alive of the women of Jabeshgilead: and yet so they sufficed them not.

datae sunt eis uxores de filiabus Iabisgalaad alias autem non

21:15 And the people repented them for Benjamin, be- reppererunt quas simili modo traderent 21:15 uniuersusque

cause that the LORD had made a breach in the tribes of Israel.

Israhel ualde doluit et egit paenitudinem super interfectione

Liber Iudicum

unius tribus ex Israhel 21:16 dixeruntque maiores natu quid 21:16 Then the elders of the congregation said, How shall

we do for wives for them that remain, seeing the women are destroyed out of Benjamin?

faciemus reliquis qui non acceperunt uxores omnes in Beniamin feminae conciderunt 21:17 et magna nobis cura in- 21:17 And they said, There must be an inheritance for

them that be escaped of Benjamin, that a tribe be not destroyed out of Israel.

gentique studio prouidendum est ne una tribus deleatur ex Israhel 21:18 filias nostras eis dare non possumus constricti 21:18 Howbeit we may not give them wives of our daugh-

ters: for the children of Israel have sworn, saying, Cursed be he that giveth a wife to Benjamin.

iuramento et maledictione qua diximus maledictus qui dederit de filiabus suis uxorem Beniamin 21:19 ceperuntque 21:19 Then they said, Behold, there is a feast of the LORD

in Shiloh yearly in a place which is on the north side of Bethel, on the east side of the highway that goeth up from

consilium atque dixerunt ecce sollemnitas Domini est in Silo Bethel to Shechem, and on the south of Lebonah. anniuersaria quae sita est ad septentrionem urbis Bethel et ad orientalem plagam uiae quae de Bethel tendit ad Sycimam et ad meridiem oppidi Lebona 21:20 praeceperuntque filiis 21:20 Therefore they commanded the children of Ben-

jamin, saying, Go and lie in wait in the vineyards;

Beniamin atque dixerunt ite et latete in uineis 21:21 cum- 21:21 And see, and, behold, if the daughters of Shiloh

come out to dance in dances, then come ye out of the vineyards, and catch you every man his wife of the daughters

que uideritis filias Silo ad ducendos choros ex more proce- of Shiloh, and go to the land of Benjamin. dere exite repente de uineis et rapite eas singuli uxores singulas et pergite in terram Beniamin 21:22 cumque uenerint 21:22 And it shall be, when their fathers or their brethren

come unto us to complain, that we will say unto them, Be favourable unto them for our sakes: because we reserved not to each man his wife in the war: for ye did not give unto them at this time, that ye should be guilty.

patres earum ac fratres et aduersum uos queri coeperint atque iurgari dicemus eis miseremini eorum non enim rapuerunt eas iure bellantium atque uictorum sed rogantibus ut acciperent non dedistis et a uestra parte peccatum est 21:23 21:23 And the children of Benjamin did so, and took them

wives, according to their number, of them that danced, whom they caught: and they went and returned unto their

feceruntque filii Beniamin ut sibi fuerat imperatum et iuxta inheritance, and repaired the cities, and dwelt in them. numerum suum rapuerunt sibi de his quae ducebant choros uxores singulas abieruntque in possessionem suam aedificantes urbes et habitantes in eis 21:24 filii quoque Israhel 21:24 And the children of Israel departed thence at that

time, every man to his tribe and to his family, and they went out from thence every man to his inheritance.

reuersi sunt per tribus et familias in tabernacula sua in diebus illis non erat rex in Israhel sed unusquisque quod sibi rectum uidebatur hoc faciebat

Biblia Sacra Vulgata

CDXLII

\underline{Liber Ruth}

1:1 in diebus unius iudicis quando iudices praeerant facta 1:1 Now it came to pass in the days when the judges ruled,

that there was a famine in the land. And a certain man of Bethlehemjudah went to sojourn in the country of Moab,

est fames in terra abiitque homo de Bethleem Iuda ut pereg- he, and his wife, and his two sons. rinaretur in regione moabitide cum uxore sua ac duobus liberis 1:2 ipse uocabatur Helimelech uxor eius Noemi e duo- 1:2 And the name of the man was Elimelech, and the name

of his wife Naomi, and the name of his two sons Mahlon and Chilion, Ephrathites of Bethlehemjudah. And they

bus filiis alter Maalon et alter Chellion Ephrathei de Beth- came into the country of Moab, and continued there. leem Iuda ingressique regionem moabitidem morabantur ibi 1:3 et mortuus est Helimelech maritus Noemi remansitque 1:3 And Elimelech Naomi’s husband died; and she was

left, and her two sons.

ipsa cum filiis 1:4 qui acceperunt uxores moabitidas qua- 1:4 And they took them wives of the women of Moab; the

name of the one was Orpah, and the name of the other Ruth: and they dwelled there about ten years.

rum una uocabatur Orpha altera Ruth manseruntque ibi decem annis 1:5 et ambo mortui sunt Maalon uidelicet et 1:5 And Mahlon and Chilion died also both of them; and

the woman was left of her two sons and her husband.

Chellion remansitque mulier orbata duobus liberis ac marito 1:6 et surrexit ut in patriam pergeret cum utraque nuru 1:6 Then she arose with her daughters in law, that she

might return from the country of Moab: for she had heard in the country of Moab how that the LORD had visited his

sua de regione moabitide audierat enim quod respexisset Do- people in giving them bread. minus populum suum et dedisset eis escas 1:7 egressa est 1:7 Wherefore she went forth out of the place where she

was, and her two daughters in law with her; and they went on the way to return unto the land of Judah.

itaque de loco peregrinationis suae cum utraque nuru et iam in uia posita reuertendi in terram Iuda 1:8 dixit ad eas ite 1:8 And Naomi said unto her two daughters in law, Go,

return each to her mother’s house: the LORD deal kindly with you, as ye have dealt with the dead, and with me.

in domum matris uestrae faciat Dominus uobiscum misericordiam sicut fecistis cum mortuis et mecum 1:9 det uobis 1:9 The LORD grant you that ye may find rest, each of

you in the house of her husband. Then she kissed them; and they lifted up their voice, and wept.

inuenire requiem in domibus uirorum quos sortiturae estis et osculata est eas quae eleuata uoce flere coeperunt 1:10 et 1:10 And they said unto her, Surely we will return with

thee unto thy people.

dicere tecum pergemus ad populum tuum 1:11 quibus illa 1:11 And Naomi said, Turn again, my daughters: why will

ye go with me? are there yet any more sons in my womb, that they may be your husbands?

respondit reuertimini filiae mi cur uenitis mecum num ultra habeo filios in utero meo ut uiros ex me sperare possitis 1:12 reuertimini filiae mi abite iam enim senectute confecta 1:12 Turn again, my daughters, go your way; for I am too

old to have an husband. If I should say, I have hope, if I should have an husband also to night, and should also bear

sum nec apta uinculo coniugali etiam si possem hac nocte sons; concipere et parere filios 1:13 si eos expectare uelitis do- 1:13 Would ye tarry for them till they were grown? would

ye stay for them from having husbands? nay, my daughters; for it grieveth me much for your sakes that the hand

nec crescant et annos impleant pubertatis ante eritis uetulae of the LORD is gone out against me. quam nubatis nolite quaeso filiae mi quia uestra angustia me magis premit et egressa est manus Domini contra me 1:14 1:14 And they lifted up their voice, and wept again: and

Orpah kissed her mother in law; but Ruth clave unto her.

eleuata igitur uoce rursum flere coeperunt Orpha osculata socrum est ac reuersa Ruth adhesit socrui suae 1:15 cui dixit 1:15 And she said, Behold, thy sister in law is gone back

unto her people, and unto her gods: return thou after thy sister in law.

Noemi en reuersa est cognata tua ad populum suum et ad deos suos uade cum ea 1:16 quae respondit ne aduerseris 1:16 And Ruth said, Intreat me not to leave thee, or to

return from following after thee: for whither thou goest, I will go; and where thou lodgest, I will lodge: thy people shall be my people, and thy God my God:

Biblia Sacra Vulgata

mihi ut relinquam te et abeam quocumque perrexeris pergam ubi morata fueris et ego pariter morabor populus tuus popu-

1:17 Where thou diest, will I die, and there will I be buried: lus meus et Deus tuus Deus meus 1:17 quae te morientem the LORD do so to me, and more also, if ought but death part thee and me.

terra susceperit in ea moriar ibique locum accipiam sepulturae haec mihi faciat Deus et haec addat si non sola mors

1:18 When she saw that she was stedfastly minded to go me et te separauerit 1:18 uidens ergo Noemi quod obstinato

with her, then she left speaking unto her.

Ruth animo decreuisset secum pergere aduersari noluit nec

1:19 So they two went until they came to Bethlehem. And ultra ad suos rediium persuadere 1:19 profectaeque sunt si-

it came to pass, when they were come to Bethlehem, that all the city was moved about them, and they said, Is this

Naomi? mul et uenerunt in Bethleem quibus urbem ingressis uelox

apud cunctos fama percrebuit dicebantque mulieres haec est

1:20 And she said unto them, Call me not Naomi, call me illa Noemi 1:20 quibus ait ne uocetis me Noemi id est pul-

Mara: for the Almighty hath dealt very bitterly with me.

chram sed uocate me Mara hoc est amaram quia ualde me

1:21 I went out full and the LORD hath brought me home amaritudine repleuit Omnipotens 1:21 egressa sum plena

again empty: why then call ye me Naomi, seeing the LORD hath testified against me, and the Almighty hath

afflicted me? et uacuam reduxit me Dominus cur igitur uocatis me Noemi 1:22 So Naomi returned, and Ruth the Moabitess, her quam humiliauit Dominus et adflixit Omnipotens 1:22 ue-

daughter in law, with her, which returned out of the country of Moab: and they came to Bethlehem in the beginning

of barley harvest. nit ergo Noemi cum Ruth Moabitide nuru sua de terra peregrinationis suae ac reuersa est in Bethleem quando primum hordea metebantur

2:1 And Naomi had a kinsman of her husband’s, a mighty 2:1 erat autem uir Helimelech consanguineus homo potens

man of wealth, of the family of Elimelech; and his name was Boaz.

et magnarum opum nomine Booz 2:2 dixitque Ruth Moa-

2:2 And Ruth the Moabitess said unto Naomi, Let me now go to the field, and glean ears of corn after him in whose sight I shall find grace. And she said unto her, Go, my daughter.

bitis ad socrum suam si iubes uadam in agrum et colligam spicas quae metentium fugerint manus ubicumque clementis in me patris familias repperero gratiam cui illa respondit

2:3 And she went, and came, and gleaned in the field after uade filia mi 2:3 abiit itaque et colligebat spicas post terga

the reapers: and her hap was to light on a part of the field belonging unto Boaz, who was of the kindred of Elim-

elech. metentium accidit autem ut ager ille haberet dominum Booz 2:4 And, behold, Boaz came from Bethlehem, and said qui erat de cognatione Helimelech 2:4 et ecce ipse ueniebat

unto the reapers, The LORD be with you. And they answered him, The LORD bless thee.

de Bethleem dixitque messoribus Dominus uobiscum qui re-

2:5 Then said Boaz unto his servant that was set over the sponderunt ei benedicat tibi Dominus 2:5 dixitque Booz iu-

reapers, Whose damsel is this?

2:6 And the servant that was set over the reapers answered ueni qui messoribus praeerat cuius est haec puella 2:6 qui

and said, It is the Moabitish damsel that came back with Naomi out of the country of Moab:

respondit haec est Moabitis quae uenit cum Noemi de re-

2:7 And she said, I pray you, let me glean and gather after gione moabitide 2:7 et rogauit ut spicas colligeret rema-

the reapers among the sheaves: so she came, and hath continued even from the morning until now, that she tarried a

little in the house. nentes sequens messorum uestigia et de mane usque nunc stat in agro et ne ad momentum quidem domum reuersa est 2:8 et ait Booz ad Ruth audi filia ne uadas ad colligendum 2:8 Then said Boaz unto Ruth, Hearest thou not, my

daughter? Go not to glean in another field, neither go from hence, but abide here fast by my maidens:

in alterum agrum nec recedas ab hoc loco sed iungere puellis meis 2:9 et ubi messuerint sequere mandaui enim pue- 2:9 Let thine eyes be on the field that they do reap, and

go thou after them: have I not charged the young men that they shall not touch thee? and when thou art athirst, go unto the vessels, and drink of that which the young men have drawn.

ris meis ut nemo tibi molestus sit sed etiam si sitieris uade ad sarcinulas et bibe aquas de quibus et pueri bibunt 2:10 2:10 Then she fell on her face, and bowed herself to the

ground, and said unto him, Why have I found grace in thine eyes, that thou shouldest take knowledge of me, see-

quae cadens in faciem suam et adorans super terram dixit ad ing I am a stranger? eum unde mihi hoc ut inuenirem gratiam ante oculos tuos et nosse me dignareris peregrinam mulierem 2:11 cui ille 2:11 And Boaz answered and said unto her, It hath fully

been shewed me, all that thou hast done unto thy mother in law since the death of thine husband: and how thou hast left thy father and thy mother, and the land of thy nativity, and art come unto a people which thou knewest not heretofore.

respondit nuntiata sunt mihi omnia quae feceris socrui tuae post mortem uiri tui et quod dereliqueris parentes tuos et terram in qua nata es et ueneris ad populum quem ante nesciebas 2:12 reddat tibi Dominus pro opere tuo et plenam 2:12 The LORD recompense thy work, and a full reward

be given thee of the LORD God of Israel, under whose wings thou art come to trust.

mercedem recipias a Domino Deo Israhel ad quem uenisti et sub cuius confugisti alas 2:13 quae ait inueni gratiam ante 2:13 Then she said, Let me find favour in thy sight, my

lord; for that thou hast comforted me, and for that thou hast spoken friendly unto thine handmaid, though I be not

oculos tuos domine mi qui consolatus es me et locutus es like unto one of thine handmaidens. ad cor ancillae tuae quae non sum similis unius puellarum tuarum 2:14 dixitque ad eam Booz quando hora uescendi 2:14 And Boaz said unto her, At mealtime come thou

hither, and eat of the bread, and dip thy morsel in the vinegar. And she sat beside the reapers: and he reached her

fuerit ueni huc et comede panem et intingue buccellam tuam parched corn, and she did eat, and was sufficed, and left. in aceto sedit itaque ad messorum latus et congessit pulentam sibi comeditque et saturata est et tulit reliquias 2:15 2:15 And when she was risen up to glean, Boaz com-

manded his young men, saying, Let her glean even among the sheaves, and reproach her not:

atque inde surrexit ut spicas ex more colligeret praecepit autem Booz pueris suis dicens etiam si uobiscum metere uoluerit ne prohibeatis eam 2:16 et de uestris quoque manipulis 2:16 And let fall also some of the handfuls of purpose for

her, and leave them, that she may glean them, and rebuke her not.

proicite de industria et remanere permittite ut absque rubore colligat et colligentem nemo corripiat 2:17 collegit ergo in 2:17 So she gleaned in the field until even, and beat out

that she had gleaned: and it was about an ephah of barley.

agro usque ad uesperam et quae collegerat uirga caedens et excutiens inuenit hordei quasi oephi mensuram id est tres modios 2:18 quos portans reuersa est in ciuitatem et osten- 2:18 And she took it up, and went into the city: and her

mother in law saw what she had gleaned: and she brought forth, and gave to her that she had reserved after she was

dit socrui suae insuper protulit et dedit ei de reliquiis cibi sui sufficed. quo saturata fuerat 2:19 dixitque ei socrus ubi hodie col- 2:19 And her mother in law said unto her, Where hast thou

gleaned to day? and where wroughtest thou? blessed be he that did take knowledge of thee. And she shewed her mother in law with whom she had wrought, and said, The man’s name with whom I wrought to day is Boaz.

legisti et ubi fecisti opus sit benedictus qui misertus est tui indicauitque ei apud quem esset operata et nomen dixit uiri quod Booz uocaretur 2:20 cui respondit Noemi benedictus 2:20 And Naomi said unto her daughter in law, Blessed be

he of the LORD, who hath not left off his kindness to the living and to the dead. And Naomi said unto her, The man is near of kin unto us, one of our next kinsmen.

Biblia Sacra Vulgata

sit a Domino quoniam eandem gratiam quam praebuerat uiuis seruauit et mortuis rursumque propinquus ait noster est

2:21 And Ruth the Moabitess said, He said unto me also, homo 2:21 et Ruth hoc quoque inquit praecepit mihi ut tam-

Thou shalt keep fast by my young men, until they have ended all my harvest.

diu messoribus eius iungerer donec omnes segetes meteren-

2:22 And Naomi said unto Ruth her daughter in law, It is tur 2:22 cui dixit socrus melius est filia mi ut cum puellis

good, my daughter, that thou go out with his maidens, that they meet thee not in any other field.

eius exeas ad metendum ne in alieno agro quispiam resistat

2:23 So she kept fast by the maidens of Boaz to glean unto tibi 2:23 iuncta est itaque puellis Booz et tamdiu cum eis

the end of barley harvest and of wheat harvest; and dwelt with her mother in law.

messuit donec hordea et triticum in horreis conderentur

3:1 Then Naomi her mother in law said unto her, My 3:1 postquam autem reuersa est ad socrum suam audiuit

daughter, shall I not seek rest for thee, that it may be well with thee?

ab ea filia mi quaeram tibi requiem et prouidebo ut bene sit

3:2 And now is not Boaz of our kindred, with whose maid- tibi 3:2 Booz iste cuius puellis in agro iuncta es propinquus

ens thou wast? Behold, he winnoweth barley to night in the threshingfloor.

est noster et hac nocte aream hordei uentilat 3:3 laua igi-

3:3 Wash thyself therefore, and anoint thee, and put thy raiment upon thee, and get thee down to the floor: but make not thyself known unto the man, until he shall have done eating and drinking.

tur et unguere et induere cultioribus uestimentis ac descende in aream non te uideat homo donec esum potumque finie-

3:4 And it shall be, when he lieth down, that thou shalt rit 3:4 quando autem ierit ad dormiendum nota locum in mark the place where he shall lie, and thou shalt go in, and uncover his feet, and lay thee down; and he will tell

thee what thou shalt do. quo dormiat ueniesque et discoperies pallium quo operitur

a parte pedum et proicies te et ibi iacebis ipse autem dicet

3:5 And she said unto her, All that thou sayest unto me I tibi quid agere debeas 3:5 quae respondit quicquid praece-

will do.

3:6 And she went down unto the floor, and did according peris faciam 3:6 descenditque in aream et fecit omnia quae

to all that her mother in law bade her.

3:7 And when Boaz had eaten and drunk, and his heart sibi imperauerat socrus 3:7 cumque comedisset Booz et bi-

was merry, he went to lie down at the end of the heap of corn: and she came softly, and uncovered his feet, and laid

her down. bisset et factus esset hilarior issetque ad dormiendum iuxta

aceruum manipulorum uenit abscondite et discoperto a pedi-

3:8 And it came to pass at midnight, that the man was bus eius pallio se proiecit 3:8 et ecce nocte iam media expa-

afraid, and turned himself: and, behold, a woman lay at his feet.

uit homo et conturbatus est uiditque mulierem iacentem ad

3:9 And he said, Who art thou? And she answered, I am pedes suos 3:9 et ait illi quae es illaque respondit ego sum

Ruth thine handmaid: spread therefore thy skirt over thine handmaid; for thou art a near kinsman.

Ruth ancilla tua expande pallium tuum super famulam tuam

3:10 And he said, Blessed be thou of the LORD, my quia propinquus es 3:10 et ille benedicta inquit es Domino

daughter: for thou hast shewed more kindness in the latter end than at the beginning, inasmuch as thou followedst not

young men, whether poor or rich. filia et priorem misericordiam posteriore superasti quia non 3:11 And now, my daughter, fear not; I will do to thee all es secuta iuuenes pauperes siue diuites 3:11 noli ergo me-

that thou requirest: for all the city of my people doth know that thou art a virtuous woman.

tuere sed quicquid dixeris mihi faciam tibi scit enim omnis populus qui habitat intra portas urbis meae mulierem te esse

3:12 And now it is true that I am thy near kinsman: how- uirtutis 3:12 nec abnuo me propinquum sed est alius me

beit there is a kinsman nearer than I. 3:13 Tarry this night, and it shall be in the morning, that if propinquior 3:13 quiesce hac nocte et facto mane si te uohe will perform unto thee the part of a kinsman, well; let him do the kinsman’s part: but if he will not do the part of a kinsman to thee, then will I do the part of a kinsman to thee, as the LORD liveth: lie down until the morning.

luerit propinquitatis iure retinere bene res acta est sin autem ille noluerit ego te absque ulla dubitatione suscipiam uiuit Dominus dormi usque mane 3:14 dormiuit itaque ad pedes 3:14 And she lay at his feet until the morning: and she

rose up before one could know another. And he said, Let it not be known that a woman came into the floor.

eius usque ad noctis abscessum surrexitque antequam homines se cognoscerent mutuo et dixit Booz caue ne quis nouerit quod huc ueneris 3:15 et rursum expande inquit palliolum 3:15 Also he said, Bring the vail that thou hast upon thee,

and hold it. And when she held it, he measured six measures of barley, and laid it on her: and she went into the

tuum quo operiris et tene utraque manu qua extendente et te- city. nente mensus est sex modios hordei et posuit super eam quae portans ingressa est ciuitatem 3:16 et uenit ad socrum suam 3:16 And when she came to her mother in law, she said,

Who art thou, my daughter? And she told her all that the man had done to her.

quae dixit ei quid egisti filia narrauitque ei omnia quae sibi fecisset homo 3:17 et ait ecce sex modios hordei dedit mihi 3:17 And she said, These six measures of barley gave he

me; for he said to me, Go not empty unto thy mother in law.

et ait nolo uacuam te reuerti ad socrum tuam 3:18 dixitque

3:18 Then said she, Sit still, my daughter, until thou know how the matter will fall: for the man will not be in rest, until he have finished the thing this day.

Noemi expecta filia donec uideamus quem res exitum habeat neque enim cessabit homo nisi conpleuerit quod locutus est

4:1 ascendit ergo Booz ad portam et sedit ibi cumque ui- 4:1 Then went Boaz up to the gate, and sat him down there:

and, behold, the kinsman of whom Boaz spake came by; unto whom he said, Ho, such a one! turn aside, sit down

disset propinquum praeterire de quo prius sermo habitus est here. And he turned aside, and sat down. dixit ad eum declina paulisper et sede hic uocans eum nomine suo qui deuertit et sedit 4:2 tollens autem Booz de- 4:2 And he took ten men of the elders of the city, and said,

Sit ye down here. And they sat down.

cem uiros de senioribus ciuitatis dixit ad eos sedete hic 4:3 4:3 And he said unto the kinsman, Naomi, that is come

again out of the country of Moab, selleth a parcel of land, which was our brother Elimelech’s:

quibus residentibus locutus est ad propinquum partem agri fratris nostri Helimelech uendit Noemi quae reuersa est de regione moabitide 4:4 quod audire te uolui et tibi dicere 4:4 And I thought to advertise thee, saying, Buy it before

the inhabitants, and before the elders of my people. If thou wilt redeem it, redeem it: but if thou wilt not redeem it, then tell me, that I may know: for there is none to redeem it beside thee; and I am after thee. And he said, I will redeem it.

coram cunctis sedentibus et maioribus natu de populo meo si uis possidere iure propinquitatis eme et posside sin autem tibi displicet hoc ipsum indica mihi ut sciam quid facere debeam nullus est enim propinquus excepto te qui prior es et me qui secundus sum at ille respondit ego agrum emam 4:5 4:5 Then said Boaz, What day thou buyest the field of

the hand of Naomi, thou must buy it also of Ruth the Moabitess, the wife of the dead, to raise up the name of

cui dixit Booz quando emeris agrum de manu mulieris Ruth the dead upon his inheritance. quoque Moabitidem quae uxor defuncti fuit debes accipere ut suscites nomen propinqui tui in hereditate sua 4:6 qui re- 4:6 And the kinsman said, I cannot redeem it for myself,

lest I mar mine own inheritance: redeem thou my right to thyself; for I cannot redeem it.

spondit cedo iure propinquitatis neque enim posteritatem familiae meae delere debeo tu meo utere priuilegio quo me libenter carere profiteor 4:7 hic autem erat mos antiquitus in 4:7 Now this was the manner in former time in Israel con-

cerning redeeming and concerning changing, for to confirm all things; a man plucked off his shoe, and gave it to his neighbour: and this was a testimony in Israel.

Biblia Sacra Vulgata

Israhel inter propinquos et si quando alter alteri suo iure cedebat ut esset firma concessio soluebat homo calciamentum suum et dabat proximo suo hoc erat testimonium cessionis in

4:8 Therefore the kinsman said unto Boaz, Buy it for thee. Israhel 4:8 dixit ergo propinquus Booz tolle calciamentum

So he drew off his shoe.

4:9 And Boaz said unto the elders, and unto all the people, quod statim soluit de pede suo 4:9 at ille maioribus natu et

Ye are witnesses this day, that I have bought all that was Elimelech’s, and all that was Chilion’s and Mahlon’s, of

the hand of Naomi. uniuerso populo testes inquit uos estis hodie quod possede-

rim omnia quae fuerunt Helimelech et Chellion et Maalon

4:10 Moreover Ruth the Moabitess, the wife of Mahlon, tradente Noemi 4:10 et Ruth Moabitidem uxorem Maalon have I purchased to be my wife, to raise up the name of the dead upon his inheritance, that the name of the dead be not cut off from among his brethren, and from the gate of his place: ye are witnesses this day.

in coniugium sumpserim ut suscitem nomen defuncti in hereditate sua ne uocabulum eius de familia sua ac fratribus et

4:11 And all the people that were in the gate, and the el- populo deleatur uos inquam huius rei testes estis 4:11 res-

ders, said, We are witnesses. The LORD make the woman that is come into thine house like Rachel and like Leah, which two did build the house of Israel: and do thou worthily in Ephratah, and be famous in Bethlehem:

pondit omnis populus qui erat in porta et maiores natu nos testes sumus faciat Dominus hanc mulierem quae ingreditur domum tuam sicut Rachel et Liam quae aedificauerunt domum Israhel ut sit exemplum uirtutis in Ephrata et habeat

4:12 And let thy house be like the house of Pharez, whom celebre nomen in Bethleem 4:12 fiatque domus tua sicut Tamar bare unto Judah, of the seed which the LORD shall give thee of this young woman.

domus Phares quem Thamar peperit Iudae de semine quod

4:13 So Boaz took Ruth, and she was his wife: and when dederit Dominus tibi ex hac puella 4:13 tulit itaque Booz

he went in unto her, the LORD gave her conception, and she bare a son.

Ruth et accepit uxorem ingressusque est ad eam et dedit illi

4:14 And the women said unto Naomi, Blessed be the Dominus ut conciperet et pareret filium 4:14 dixeruntque

LORD, which hath not left thee this day without a kinsman, that his name may be famous in Israel.

mulieres ad Noemi benedictus Dominus qui non est passus ut deficeret successor familiae tuae et uocaretur nomen eius

4:15 And he shall be unto thee a restorer of thy life, and a in Israhel 4:15 et habeas qui consoletur animam tuam et nourisher of thine old age: for thy daughter in law, which loveth thee, which is better to thee than seven sons, hath

born him. enutriat senectutem de nuru enim tua natus est quae te diligit 4:16 And Naomi took the child, and laid it in her bosom, et multo tibi est melior quam si septem haberes filios 4:16

and became nurse unto it.

susceptumque Noemi puerum posuit in sinu suo et nutricis

4:17 And the women her neighbours gave it a name, say- ac gerulae officio fungebatur 4:17 uicinae autem mulieres

ing, There is a son born to Naomi; and they called his name Obed: he is the father of Jesse, the father of David.

congratulantes ei et dicentes natus est filius Noemi uocaue-

4:18 Now these are the generations of Pharez: Pharez be- runt nomen eius Obed hic est pater Isai patris Dauid 4:18 gat Hezron,

4:19 And Hezron begat Ram, and Ram begat Amminadab, hae sunt generationes Phares Phares genuit Esrom 4:19 Es-

4:20 And Amminadab begat Nahshon, and Nahshon begat rom genuit Aram Aram genuit Aminadab 4:20 Aminadab Salmon,

4:21 And Salmon begat Boaz, and Boaz begat Obed, genuit Naasson Naasson genuit Salma 4:21 Salma genuit 4:22 And Obed begat Jesse, and Jesse begat David. Booz Booz genuit Obed 4:22 Obed genuit Isai Isai genuit

\underline{Liber Ruth}

Dauid

Biblia Sacra Vulgata

CDL

Liber I Samuelis

1:1 fuit uir unus de Ramathaimsophim de monte Ephraim 1:1 Now there was a certain man of Ramathaimzophim,

of mount Ephraim, and his name was Elkanah, the son of Jeroham, the son of Elihu, the son of Tohu, the son of

et nomen eius Helcana filius Hieroam filii Heliu filii Thau Zuph, an Ephrathite: filii Suph Ephratheus 1:2 et habuit duas uxores nomen uni 1:2 And he had two wives; the name of the one was Han-

nah, and the name of the other Peninnah: and Peninnah had children, but Hannah had no children.

Anna et nomen secundae Fenenna fueruntque Fenennae filii Annae autem non erant liberi 1:3 et ascendebat uir ille 1:3 And this man went up out of his city yearly to worship

and to sacrifice unto the LORD of hosts in Shiloh. And the two sons of Eli, Hophni and Phinehas, the priests of

de ciuitate sua statutis diebus ut adoraret et sacrificaret Do- the LORD, were there. mino exercituum in Silo erant autem ibi duo filii Heli Ofni et Finees sacerdotes Domini 1:4 uenit ergo dies et immolauit 1:4 And when the time was that Elkanah offered, he gave

to Peninnah his wife, and to all her sons and her daughters, portions:

Helcana deditque Fenennae uxori suae et cunctis filiis eius et filiabus partes 1:5 Annae autem dedit partem unam tristis 1:5 But unto Hannah he gave a worthy portion; for he

loved Hannah: but the LORD had shut up her womb.

quia Annam diligebat Dominus autem concluserat uuluam

eius 1:6 adfligebat quoque eam aemula eius et uehemen- 1:6 And her adversary also provoked her sore, for to make

her fret, because the LORD had shut up her womb.

ter angebat in tantum ut exprobraret quod conclusisset Dominus uuluam eius 1:7 sicque faciebat per singulos annos 1:7 And as he did so year by year, when she went up to the

house of the LORD, so she provoked her; therefore she wept, and did not eat.

cum redeente tempore ascenderent templum Domini et sic prouocabat eam porro illa flebat et non capiebat cibum 1:8 1:8 Then said Elkanah her husband to her, Hannah, why

weepest thou? and why eatest thou not? and why is thy heart grieved? am not I better to thee than ten sons?

dixit ergo ei Helcana uir suus Anna cur fles et quare non comedis et quam ob rem adfligitur cor tuum numquid non ego melior sum tibi quam decem filii 1:9 surrexit autem Anna 1:9 So Hannah rose up after they had eaten in Shiloh, and

after they had drunk. Now Eli the priest sat upon a seat by a post of the temple of the LORD.

postquam comederat in Silo et biberat et Heli sacerdote sedente super sellam ante postes templi Domini 1:10 cum es- 1:10 And she was in bitterness of soul, and prayed unto

the LORD, and wept sore.

set amaro animo orauit Dominum flens largiter 1:11 et uo- 1:11 And she vowed a vow, and said, O LORD of hosts, if

thou wilt indeed look on the affliction of thine handmaid, and remember me, and not forget thine handmaid, but wilt give unto thine handmaid a man child, then I will give him unto the LORD all the days of his life, and there shall no razor come upon his head.

tum uouit dicens Domine exercituum si respiciens uideris adflictionem famulae tuae et recordatus mei fueris nec oblitus ancillae tuae dederisque seruae tuae sexum uirilem dabo eum Domino omnes dies uitae eius et nouacula non ascendet super caput eius 1:12 factum est ergo cum illa multiplica- 1:12 And it came to pass, as she continued praying before

the LORD, that Eli marked her mouth.

ret preces coram Domino ut Heli obseruaret os eius 1:13 1:13 Now Hannah, she spake in her heart; only her lips

moved, but her voice was not heard: therefore Eli thought she had been drunken.

porro Anna loquebatur in corde suo tantumque labia illius mouebantur et uox penitus non audiebatur aestimauit igitur eam Heli temulentam 1:14 dixitque ei usquequo ebria eris 1:14 And Eli said unto her, How long wilt thou be

drunken? put away thy wine from thee.

digere paulisper uinum quo mades 1:15 respondens Anna 1:15 And Hannah answered and said, No, my lord, I am

a woman of a sorrowful spirit: I have drunk neither wine nor strong drink, but have poured out my soul before the

nequaquam inquit domine mi nam mulier infelix nimis ego LORD.

Biblia Sacra Vulgata

sum uinumque et omne quod inebriare potest non bibi sed

1:16 Count not thine handmaid for a daughter of Belial: effudi animam meam in conspectu Domini 1:16 ne reputes for out of the abundance of my complaint and grief have I spoken hitherto.

ancillam tuam quasi unam de filiabus Belial quia ex multitudine doloris et maeroris mei locuta sum usque in praesens

1:17 Then Eli answered and said, Go in peace: and the 1:17 tunc Heli ait ei uade in pace et Deus Israhel det tibi peti-

God of Israel grant thee thy petition that thou hast asked of him.

tionem quam rogasti eum 1:18 et illa dixit utinam inueniat

1:18 And she said, Let thine handmaid find grace in thy sight. So the woman went her way, and did eat, and her countenance was no more sad.

ancilla tua gratiam in oculis tuis et abiit mulier in uiam suam et comedit uultusque eius non sunt amplius in diuersa mu-

1:19 And they rose up in the morning early, and wor- tati 1:19 et surrexerunt mane et adorauerunt coram Domino

shipped before the LORD, and returned, and came to their house to Ramah: and Elkanah knew Hannah his wife; and

the LORD remembered her. reuersique sunt et uenerunt in domum suam Ramatha cog-

nouit autem Helcana Annam uxorem suam et recordatus est

1:20 Wherefore it came to pass, when the time was come eius Dominus 1:20 et factum est post circulum dierum con-

about after Hannah had conceived, that she bare a son, and called his name Samuel, saying, Because I have asked him

of the LORD. cepit Anna et peperit filium uocauitque nomen eius Samuhel 1:21 And the man Elkanah, and all his house, went up to eo quod a Domino postulasset eum 1:21 ascendit autem uir

offer unto the LORD the yearly sacrifice, and his vow.

Helcana et omnis domus eius ut immolaret Domino hostiam

1:22 But Hannah went not up; for she said unto her hus- sollemnem et uotum suum 1:22 et Anna non ascendit dixit

band, I will not go up until the child be weaned, and then I will bring him, that he may appear before the LORD, and

there abide for ever. enim uiro suo non uadam donec ablactetur infans et ducam

eum et appareat ante conspectum Domini et maneat ibi iu-

1:23 And Elkanah her husband said unto her, Do what giter 1:23 et ait ei Helcana uir suus fac quod bonum tibi

seemeth thee good; tarry until thou have weaned him; only the LORD establish his word. So the woman abode, and

gave her son suck until she weaned him. uidetur et mane donec ablactes eum precorque ut impleat

Dominus uerbum suum mansit ergo mulier et lactauit filium

1:24 And when she had weaned him, she took him up with suum donec amoueret eum a lacte 1:24 et adduxit eum se-

her, with three bullocks, and one ephah of flour, and a bottle of wine, and brought him unto the house of the LORD

in Shiloh: and the child was young. cum postquam ablactauerat in uitulis tribus et tribus modiis

farinae et amphora uini et adduxit eum ad domum Domini in

1:25 And they slew a bullock, and brought the child to Eli. Silo puer autem erat adhuc infantulus 1:25 et immolauerunt

1:26 And she said, Oh my lord, as thy soul liveth, my lord, uitulum et obtulerunt puerum Heli 1:26 et ait obsecro mi I am the woman that stood by thee here, praying unto the LORD.

domine uiuit anima tua domine ego sum illa mulier quae steti

1:27 For this child I prayed; and the LORD hath given me coram te hic orans Dominum 1:27 pro puero isto oraui et my petition which I asked of him:

dedit Dominus mihi petitionem meam quam postulaui eum

1:28 Therefore also I have lent him to the LORD; as long 1:28 idcirco et ego commodaui eum Domino cunctis diebus

as he liveth he shall be lent to the LORD. And he worshipped the LORD there.

quibus fuerit accommodatus Domino et adorauerunt ibi Dominum et orauit Anna et ait

2:1 And Hannah prayed, and said, My heart rejoiceth in 2:1 exultauit cor meum in Domino exaltatum est cornu the LORD, mine horn is exalted in the LORD: my mouth is enlarged over mine enemies; because I rejoice in thy salvation.

Liber I Samuelis

meum in Domino dilatatum est os meum super inimicos meos quia laetata sum in salutari tuo 2:2 non est sanctus 2:2 There is none holy as the LORD: for there is none

beside thee: neither is there any rock like our God.

ut est Dominus neque enim est alius extra te et non est fortis sicut Deus noster 2:3 nolite multiplicare loqui sublimia 2:3 Talk no more so exceeding proudly; let not arrogancy

come out of your mouth: for the LORD is a God of knowledge, and by him actions are weighed.

gloriantes recedant uetera de ore uestro quoniam Deus scientiarum Dominus est et ipsi praeparantur cogitationes 2:4 2:4 The bows of the mighty men are broken, and they that

stumbled are girded with strength.

arcus fortium superatus est et infirmi accincti sunt robore 2:5 saturati prius pro pane se locauerunt et famelici saturati 2:5 They that were full have hired out themselves for

bread; and they that were hungry ceased: so that the barren hath born seven; and she that hath many children is

sunt donec sterilis peperit plurimos et quae multos habebat waxed feeble. filios infirmata est 2:6 Dominus mortificat et uiuificat de- 2:6 The LORD killeth, and maketh alive: he bringeth

down to the grave, and bringeth up.

ducit ad infernum et reducit 2:7 Dominus pauperem facit 2:7 The LORD maketh poor, and maketh rich: he bringeth

low, and lifteth up.

et ditat humiliat et subleuat 2:8 suscitat de puluere egenum 2:8 He raiseth up the poor out of the dust, and lifteth up the

beggar from the dunghill, to set them among princes, and to make them inherit the throne of glory: for the pillars of the earth are the LORD’s, and he hath set the world upon them.

et de stercore eleuat pauperem ut sedeat cum principibus et solium gloriae teneat Domini enim sunt cardines terrae et

posuit super eos orbem 2:9 pedes sanctorum suorum ser- 2:9 He will keep the feet of his saints, and the wicked shall

be silent in darkness; for by strength shall no man prevail.

uabit et impii in tenebris conticescent quia non in fortitudine roborabitur uir 2:10 Dominum formidabunt aduersarii eius 2:10 The adversaries of the LORD shall be broken to

pieces; out of heaven shall he thunder upon them: the LORD shall judge the ends of the earth; and he shall give

super ipsos in caelis tonabit Dominus iudicabit fines terrae strength unto his king, and exalt the horn of his anointed. et dabit imperium regi suo et sublimabit cornu christi sui 2:11 et abiit Helcana Ramatha in domum suam puer autem 2:11 And Elkanah went to Ramah to his house. And the

child did minister unto the LORD before Eli the priest.

erat minister in conspectu Domini ante faciem Heli sacerdotis 2:12 porro filii Heli filii Belial nescientes Dominum 2:12 Now the sons of Eli were sons of Belial; they knew

not the LORD.

2:13 neque officium sacerdotum ad populum sed quicumque 2:13 And the priest’s custom with the people was, that,

when any man offered sacrifice, the priest’s servant came, while the flesh was in seething, with a fleshhook of three

immolasset uictimam ueniebat puer sacerdotis dum coque- teeth in his hand; rentur carnes et habebat fuscinulam tridentem in manu sua 2:14 et mittebat eam in lebetem uel in caldariam aut in ollam 2:14 And he struck it into the pan, or kettle, or caldron, or

pot; all that the fleshhook brought up the priest took for himself. So they did in Shiloh unto all the Israelites that

siue in caccabum et omne quod leuabat fuscinula tollebat came thither. sacerdos sibi sic faciebant uniuerso Israheli uenientium in Silo 2:15 etiam antequam adolerent adipem ueniebat puer 2:15 Also before they burnt the fat, the priest’s servant

came, and said to the man that sacrificed, Give flesh to roast for the priest; for he will not have sodden flesh of

sacerdotis et dicebat immolanti da mihi carnem ut coquam thee, but raw. sacerdoti non enim accipiam a te carnem coctam sed crudam 2:16 dicebatque illi immolans incendatur primum iuxta mo- 2:16 And if any man said unto him, Let them not fail to

burn the fat presently, and then take as much as thy soul desireth; then he would answer him, Nay; but thou shalt

rem hodie adeps et tolle tibi quantumcumque desiderat an- give it me now: and if not, I will take it by force.

Biblia Sacra Vulgata

ima tua qui respondens aiebat ei nequaquam nunc enim da-

2:17 Wherefore the sin of the young men was very great bis alioquin tollam ui 2:17 erat ergo peccatum puerorum

before the LORD: for men abhorred the offering of the LORD.

grande nimis coram Domino quia detrahebant homines sac-

2:18 But Samuel ministered before the LORD, being a rificio Domini 2:18 Samuhel autem ministrabat ante faciem

child, girded with a linen ephod.

2:19 Moreover his mother made him a little coat, and Domini puer accinctus ephod lineo 2:19 et tunicam paruam

brought it to him from year to year, when she came up with her husband to offer the yearly sacrifice.

faciebat ei mater sua quam adferebat statutis diebus ascen-

2:20 And Eli blessed Elkanah and his wife, and said, The dens cum uiro suo ut immolaret hostiam sollemnem 2:20 et

LORD give thee seed of this woman for the loan which is lent to the LORD. And they went unto their own home.

benedixit Heli Helcanae et uxori eius dixitque reddat Dominus tibi semen de muliere hac pro fenore quod commodasti

2:21 And the LORD visited Hannah, so that she con- Domino et abierunt in locum suum 2:21 uisitauit ergo Do-

ceived, and bare three sons and two daughters. And the child Samuel grew before the LORD.

minus Annam et concepit et peperit tres filios et duas filias

2:22 Now Eli was very old, and heard all that his sons did et magnificatus est puer Samuhel apud Dominum 2:22 Heli unto all Israel; and how they lay with the women that assembled at the door of the tabernacle of the congregation.

autem erat senex ualde et audiuit omnia quae faciebant filii sui uniuerso Israheli et quomodo dormiebant cum mulieri-

2:23 And he said unto them, Why do ye such things? for I bus quae obseruabant ad ostium tabernaculi 2:23 et dixit

hear of your evil dealings by all this people.

eis quare facitis res huiuscemodi quas ego audio res pessi-

2:24 Nay, my sons; for it is no good report that I hear: ye mas ab omni populo 2:24 nolite filii mi non enim est bona make the LORD’s people to transgress.

fama quam ego audio ut transgredi faciatis populum Domini

2:25 If one man sin against another, the judge shall judge 2:25 si peccauerit uir in uirum placari ei potest Deus si au-

him: but if a man sin against the LORD, who shall intreat for him? Notwithstanding they hearkened not unto the voice of their father, because the LORD would slay them.

tem in Domino peccauerit uir quis orabit pro eo et non audierunt uocem patris sui quia uoluit Dominus occidere eos

2:26 And the child Samuel grew on, and was in favour 2:26 puer autem Samuhel proficiebat atque crescebat et pla-

both with the LORD, and also with men.

2:27 And there came a man of God unto Eli, and said unto cebat tam Deo quam hominibus 2:27 uenit autem uir Dei ad

him, Thus saith the LORD, Did I plainly appear unto the house of thy father, when they were in Egypt in Pharaoh’s

house? Heli et ait ad eum haec dicit Dominus numquid non aperte

reuelatus sum domui patris tui cum essent in Aegypto in

2:28 And did I choose him out of all the tribes of Israel to domo Pharaonis 2:28 et elegi eum ex omnibus tribubus Isbe my priest, to offer upon mine altar, to burn incense, to wear an ephod before me? and did I give unto the house of thy father all the offerings made by fire of the children of Israel?

rahel mihi in sacerdotem ut ascenderet altare meum et adoleret mihi incensum et portaret ephod coram me et dedi domui

2:29 Wherefore kick ye at my sacrifice and at mine of- patris tui omnia de sacrificiis filiorum Israhel 2:29 quare

fering, which I have commanded in my habitation; and honourest thy sons above me, to make yourselves fat with

the chiefest of all the offerings of Israel my people? calce abicitis uictimam meam et munera mea quae praecepi ut offerrentur in templo et magis honorasti filios tuos quam me ut comederetis primitias omnis sacrificii Israhel populi

2:30 Wherefore the LORD God of Israel saith, I said in- mei 2:30 propterea ait Dominus Deus Israhel loquens lo-

deed that thy house, and the house of thy father, should walk before me for ever: but now the LORD saith, Be it far from me; for them that honour me I will honour, and they that despise me shall be lightly esteemed.

Liber I Samuelis

cutus sum ut domus tua et domus patris tui ministraret in conspectu meo usque in sempiternum nunc autem dicit Dominus absit hoc a me sed quicumque glorificauerit me glorificabo eum qui autem contemnunt me erunt ignobiles 2:31 2:31 Behold, the days come, that I will cut off thine arm,

and the arm of thy father’s house, that there shall not be an old man in thine house.

ecce dies ueniunt et praecidam brachium tuum et brachium domus patris tui ut non sit senex in domo tua 2:32 et uide- 2:32 And thou shalt see an enemy in my habitation, in all

the wealth which God shall give Israel: and there shall not be an old man in thine house for ever.

bis aemulum tuum in templo in uniuersis prosperis Israhel et non erit senex in domo tua omnibus diebus 2:33 uerum- 2:33 And the man of thine, whom I shall not cut off from

mine altar, shall be to consume thine eyes, and to grieve thine heart: and all the increase of thine house shall die in

tamen non auferam penitus uirum ex te ab altari meo sed ut the flower of their age. deficiant oculi tui et tabescat anima tua et pars magna domus tuae morietur cum ad uirilem aetatem uenerit 2:34 hoc au- 2:34 And this shall be a sign unto thee, that shall come

upon thy two sons, on Hophni and Phinehas; in one day they shall die both of them.

tem erit tibi signum quod uenturum est duobus filiis tuis Ofni et Finees in die uno morientur ambo 2:35 et suscitabo mihi 2:35 And I will raise me up a faithful priest, that shall do

according to that which is in mine heart and in my mind: and I will build him a sure house; and he shall walk before

sacerdotem fidelem qui iuxta cor meum et animam meam mine anointed for ever. faciat et aedificabo ei domum fidelem et ambulabit coram christo meo cunctis diebus 2:36 futurum est autem ut qui- 2:36 And it shall come to pass, that every one that is left

in thine house shall come and crouch to him for a piece of silver and a morsel of bread, and shall say, Put me, I pray thee, into one of the priests’ offices, that I may eat a piece of bread.

cumque remanserit in domo tua ueniat ut oretur pro eo et offerat nummum argenteum et tortam panis dicatque dimitte me obsecro ad unam partem sacerdotalem ut comedam buccellam panis

3:1 puer autem Samuhel ministrabat Domino coram Heli 3:1 And the child Samuel ministered unto the LORD be-

fore Eli. And the word of the LORD was precious in those days; there was no open vision.

et sermo Domini erat pretiosus in diebus illis non erat uisio manifesta 3:2 factum est ergo in die quadam Heli iacebat in 3:2 And it came to pass at that time, when Eli was laid

down in his place, and his eyes began to wax dim, that he could not see;

loco suo et oculi eius caligauerant nec poterat uidere 3:3 lu-

3:3 And ere the lamp of God went out in the temple of the LORD, where the ark of God was, and Samuel was laid down to sleep;

cerna Dei antequam extingueretur Samuhel autem dormiebat in templo Domini ubi erat arca Dei 3:4 et uocauit Dominus 3:4 That the LORD called Samuel: and he answered, Here

am I.

Samuhel qui respondens ait ecce ego 3:5 et cucurrit ad Heli 3:5 And he ran unto Eli, and said, Here am I; for thou

calledst me. And he said, I called not; lie down again. And he went and lay down.

et dixit ecce ego uocasti enim me qui dixit non uocaui reuertere dormi et abiit et dormiuit 3:6 et adiecit Dominus uo- 3:6 And the LORD called yet again, Samuel. And Samuel

arose and went to Eli, and said, Here am I; for thou didst call me. And he answered, I called not, my son; lie down

care rursum Samuhel consurgensque Samuhel abiit ad Heli again. et dixit ecce ego quia uocasti me qui respondit non uocaui te fili mi reuertere et dormi 3:7 porro Samuhel necdum scie- 3:7 Now Samuel did not yet know the LORD, neither was

the word of the LORD yet revealed unto him.

bat Dominum neque reuelatus fuerat ei sermo Domini 3:8 3:8 And the LORD called Samuel again the third time.

And he arose and went to Eli, and said, Here am I; for thou didst call me. And Eli perceived that the LORD had called the child.

Biblia Sacra Vulgata

et adiecit Dominus et uocauit adhuc Samuhel tertio qui con-

3:9 Therefore Eli said unto Samuel, Go, lie down: and it surgens abiit ad Heli 3:9 et ait ecce ego quia uocasti me shall be, if he call thee, that thou shalt say, Speak, LORD; for thy servant heareth. So Samuel went and lay down in

his place. intellexit igitur Heli quia Dominus uocaret puerum et ait ad Samuhel uade et dormi et si deinceps uocauerit te dices loquere Domine quia audit seruus tuus abiit ergo Samuhel et

3:10 And the LORD came, and stood, and called as at dormiuit in loco suo 3:10 et uenit Dominus et stetit et uo-

other times, Samuel, Samuel. Then Samuel answered, Speak; for thy servant heareth.

cauit sicut uocauerat secundo Samuhel Samuhel et ait Sa-

3:11 And the LORD said to Samuel, Behold, I will do a muhel loquere quia audit seruus tuus 3:11 et dixit Domi-

thing in Israel, at which both the ears of every one that heareth it shall tingle.

nus ad Samuhel ecce ego facio uerbum in Israhel quod qui-

3:12 In that day I will perform against Eli all things which cumque audierit tinnient ambae aures eius 3:12 in die illo I have spoken concerning his house: when I begin, I will also make an end.

suscitabo aduersum Heli omnia quae locutus sum super do-

3:13 For I have told him that I will judge his house for ever mum eius incipiam et conplebo 3:13 praedixi enim ei quod for the iniquity which he knoweth; because his sons made themselves vile, and he restrained them not.

iudicaturus essem domum eius in aeternum propter iniquitatem eo quod nouerat indigne agere filios suos et non corri-

3:14 And therefore I have sworn unto the house of Eli, puit eos 3:14 idcirco iuraui domui Heli quod non expietur

that the iniquity of Eli’s house shall not be purged with sacrifice nor offering for ever.

iniquitas domus eius uictimis et muneribus usque in aeter-

3:15 And Samuel lay until the morning, and opened the num 3:15 dormiuit autem Samuhel usque mane aperuitque

doors of the house of the LORD. And Samuel feared to shew Eli the vision.

ostia domus Domini et Samuhel timebat indicare uisionem

3:16 Then Eli called Samuel, and said, Samuel, my son. Heli 3:16 uocauit ergo Heli Samuhelem et dixit Samuhel

And he answered, Here am I.

3:17 And he said, What is the thing that the LORD hath fili mi qui respondens ait praesto sum 3:17 et interrogauit

said unto thee? I pray thee hide it not from me: God do so to thee, and more also, if thou hide any thing from me of

all the things that he said unto thee. eum quis est sermo quem locutus est ad te oro te ne celaue-

ris me haec faciat tibi Deus et haec addat si absconderis a

3:18 And Samuel told him every whit, and hid nothing me sermonem ex omnibus uerbis quae dicta sunt tibi 3:18

from him. And he said, It is the LORD: let him do what seemeth him good.

indicauit itaque ei Samuhel uniuersos sermones et non abscondit ab eo et ille respondit Dominus est quod bonum est

3:19 And Samuel grew, and the LORD was with him, and in oculis suis faciat 3:19 creuit autem Samuhel et Domi-

did let none of his words fall to the ground.

nus erat cum eo et non cecidit ex omnibus uerbis eius in

3:20 And all Israel from Dan even to Beersheba knew that terram 3:20 et cognouit uniuersus Israhel a Dan usque Ber-

Samuel was established to be a prophet of the LORD.

3:21 And the LORD appeared again in Shiloh: for the sabee quod fidelis Samuhel propheta esset Domini 3:21 et

LORD revealed himself to Samuel in Shiloh by the word of the LORD.

addidit Dominus ut appareret in Silo quoniam reuelatus fuerat Dominus Samuheli in Silo iuxta uerbum Domini et euenit sermo Samuhelis uniuerso Israheli

4:1 And the word of Samuel came to all Israel. Now Is- 4:1 egressus est namque Israhel obuiam Philisthim in pro-

rael went out against the Philistines to battle, and pitched beside Ebenezer: and the Philistines pitched in Aphek.

Liber I Samuelis

elium et castrametatus est iuxta lapidem Adiutorii porro Phi-

listhim uenerunt in Afec 4:2 et instruxerunt aciem contra 4:2 And the Philistines put themselves in array against Israel: and when they joined battle, Israel was smitten before the Philistines: and they slew of the army in the field

Israhel inito autem certamine terga uertit Israhel Philistheis about four thousand men. et caesa sunt in illo certamine passim per agros quasi quattuor milia uirorum 4:3 et reuersus est populus ad castra 4:3 And when the people were come into the camp, the

elders of Israel said, Wherefore hath the LORD smitten us to day before the Philistines? Let us fetch the ark of the covenant of the LORD out of Shiloh unto us, that, when it cometh among us, it may save us out of the hand of our enemies.

dixeruntque maiores natu de Israhel quare percussit nos Dominus hodie coram Philisthim adferamus ad nos de Silo arcam foederis Domini et ueniat in medium nostri ut saluet nos de manu inimicorum nostrorum 4:4 misit ergo populus 4:4 So the people sent to Shiloh, that they might bring

from thence the ark of the covenant of the LORD of hosts, which dwelleth between the cherubims: and the two sons of Eli, Hophni and Phinehas, were there with the ark of the covenant of God.

in Silo et tulerunt inde arcam foederis Domini exercituum sedentis super cherubin erantque duo filii Heli cum arca foederis Domini Ofni et Finees 4:5 cumque uenisset arca foe- 4:5 And when the ark of the covenant of the LORD came

into the camp, all Israel shouted with a great shout, so that the earth rang again.

deris Domini in castra uociferatus est omnis Israhel clamore grandi et personuit terra 4:6 et audierunt Philisthim uocem 4:6 And when the Philistines heard the noise of the shout,

they said, What meaneth the noise of this great shout in the camp of the Hebrews? And they understood that the

clamoris dixeruntque quaenam haec est uox clamoris magni ark of the LORD was come into the camp. in castris Hebraeorum et cognouerunt quod arca Domini uenisset in castra 4:7 timueruntque Philisthim dicentes uenit 4:7 And the Philistines were afraid, for they said, God is

come into the camp. And they said, Woe unto us! for there hath not been such a thing heretofore.

Deus in castra et ingemuerunt 4:8 uae nobis non enim fuit

4:8 Woe unto us! who shall deliver us out of the hand of these mighty Gods? these are the Gods that smote the Egyptians with all the plagues in the wilderness.

tanta exultatio heri et nudius tertius uae nobis quis nos seruabit de manu deorum sublimium istorum hii sunt dii qui percusserunt Aegyptum omni plaga in deserto 4:9 confor- 4:9 Be strong and quit yourselves like men, O ye

Philistines, that ye be not servants unto the Hebrews, as they have been to you: quit yourselves like men, and fight.

tamini et estote uiri Philisthim ne seruiatis Hebraeis sicut illi seruierunt uobis confortamini et bellate 4:10 pugnauerunt 4:10 And the Philistines fought, and Israel was smitten,

and they fled every man into his tent: and there was a very great slaughter; for there fell of Israel thirty thousand foot-

ergo Philisthim et caesus est Israhel et fugit unusquisque in men. tabernaculum suum et facta est plaga magna nimis et ceciderunt de Israhel triginta milia peditum 4:11 et arca Dei capta 4:11 And the ark of God was taken; and the two sons of

Eli, Hophni and Phinehas, were slain.

est duoque filii Heli mortui sunt Ofni et Finees 4:12 currens 4:12 And there ran a man of Benjamin out of the army,

and came to Shiloh the same day with his clothes rent, and with earth upon his head.

autem uir de Beniamin ex acie uenit in Silo in die illo scissa ueste et conspersus puluere caput 4:13 cumque ille uenisset 4:13 And when he came, lo, Eli sat upon a seat by the

wayside watching: for his heart trembled for the ark of God. And when the man came into the city, and told it, all

Heli sedebat super sellam contra uiam aspectans erat enim the city cried out. cor eius pauens pro arca Domini uir autem ille postquam ingressus est nuntiauit urbi et ululauit omnis ciuitas 4:14 et 4:14 And when Eli heard the noise of the crying, he said,

What meaneth the noise of this tumult? And the man came in hastily, and told Eli.

audiuit Heli sonitum clamoris dixitque quis est hic sonitus

Biblia Sacra Vulgata

tumultus huius at ille festinauit et uenit et adnuntiauit Heli

4:15 Now Eli was ninety and eight years old; and his eyes 4:15 Heli autem erat nonaginta et octo annorum et oculi eius

were dim, that he could not see.

4:16 And the man said unto Eli, I am he that came out of caligauerant et uidere non poterat 4:16 et dixit ad Heli ego

the army, and I fled to day out of the army. And he said, What is there done, my son?

sum qui ueni de proelio et ego qui de acie fugi hodie cui ille

4:17 And the messenger answered and said, Israel is fled ait quid actum est fili mi 4:17 respondens autem qui nunbefore the Philistines, and there hath been also a great slaughter among the people, and thy two sons also, Hophni

and Phinehas, are dead, and the ark of God is taken. tiabat fugit inquit Israhel coram Philisthim et ruina magna

facta est in populo insuper et duo filii tui mortui sunt Ofni

4:18 And it came to pass, when he made mention of the ark et Finees et arca Dei capta est 4:18 cumque ille nominasset

of God, that he fell from off the seat backward by the side of the gate, and his neck brake, and he died: for he was an

old man, and heavy. And he had judged Israel forty years. arcam Dei cecidit de sella retrorsum iuxta ostium et fractis

ceruicibus mortuus est senex enim erat uir et grandeuus et

4:19 And his daughter in law, Phinehas’ wife, was with ipse iudicauit Israhel quadraginta annis 4:19 nurus autem

child, near to be delivered: and when she heard the tidings that the ark of God was taken, and that her father in law and her husband were dead, she bowed herself and travailed; for her pains came upon her.

eius uxor Finees praegnans erat uicinaque partui et audito nuntio quod capta esset arca Dei et mortuus socer suus et uir suus incuruauit se et peperit inruerant enim in eam dolores

4:20 And about the time of her death the women that stood subiti 4:20 in ipso autem momento mortis eius dixerunt ei

by her said unto her, Fear not; for thou hast born a son. But she answered not, neither did she regard it.

quae stabant circa eam ne timeas quia filium peperisti quae

4:21 And she named the child Ichabod, saying, The glory non respondit eis neque animaduertit 4:21 et uocauit pue-

is departed from Israel: because the ark of God was taken, and because of her father in law and her husband.

rum Hicabod dicens translata est gloria de Israhel quia capta

4:22 And she said, The glory is departed from Israel: for est arca Dei et pro socero suo et pro uiro suo 4:22 et ait the ark of God is taken.

translata est gloria ab Israhel eo quod capta esset arca Dei

5:1 And the Philistines took the ark of God, and brought it 5:1 Philisthim autem tulerunt arcam Dei et asportauerunt

from Ebenezer unto Ashdod. 5:2 When the Philistines took the ark of God, they brought eam a lapide Adiutorii in Azotum 5:2 tulerunt Philisthim it into the house of Dagon, and set it by Dagon.

arcam Dei et intulerunt eam in templum Dagon et statuerunt

5:3 And when they of Ashdod arose early on the morrow, eam iuxta Dagon 5:3 cumque surrexissent diluculo Azo-

behold, Dagon was fallen upon his face to the earth before the ark of the LORD. And they took Dagon, and set him

in his place again. tii altera die ecce Dagon iacebat pronus in terram ante ar-

cam Domini et tulerunt Dagon et restituerunt eum in loco

5:4 And when they arose early on the morrow morning, suo 5:4 rursumque mane die alio consurgentes inuenerunt

behold, Dagon was fallen upon his face to the ground before the ark of the LORD; and the head of Dagon and both the palms of his hands were cut off upon the threshold; only the stump of Dagon was left to him.

Dagon iacentem super faciem suam in terram coram arca Domini caput autem Dagon et duae palmae manuum eius

5:5 Therefore neither the priests of Dagon, nor any that abscisae erant super limen 5:5 porro Dagon truncus solus

come into Dagon’s house, tread on the threshold of Dagon in Ashdod unto this day.

remanserat in loco suo propter hanc causam non calcant sacerdotes Dagon et omnes qui ingrediuntur templum eius su-

5:6 But the hand of the LORD was heavy upon them of per limen Dagon in Azoto usque in hodiernum diem 5:6

Ashdod, and he destroyed them, and smote them with emerods, even Ashdod and the coasts thereof.

Liber I Samuelis

adgrauata autem est manus Domini super Azotios et demolitus est eos et percussit in secretiori parte natium Azotum et fines eius 5:7 uidentes autem uiri azotii huiuscemodi pla- 5:7 And when the men of Ashdod saw that it was so, they

said, The ark of the God of Israel shall not abide with us: for his hand is sore upon us, and upon Dagon our god.

gam dixerunt non maneat arca Dei Israhel apud nos quoniam dura est manus eius super nos et super Dagon deum nostrum 5:8 et mittentes congregauerunt omnes satrapas Philisthino- 5:8 They sent therefore and gathered all the lords of the

Philistines unto them, and said, What shall we do with the ark of the God of Israel? And they answered, Let the ark of the God of Israel be carried about unto Gath. And they carried the ark of the God of Israel about thither.

rum ad se et dixerunt quid faciemus de arca Dei Israhel responderuntque Getthei circumducatur arca Dei Israhel et cir-

cumduxerunt arcam Dei Israhel 5:9 illis autem circumdu- 5:9 And it was so, that, after they had carried it about, the hand of the LORD was against the city with a very great destruction: and he smote the men of the city, both small

centibus eam fiebat manus Dei per singulas ciuitates inter- and great, and they had emerods in their secret parts. fectionis magnae nimis et percutiebat uiros uniuscuiusque urbis a paruo usque ad maiorem et conputrescebant prominentes extales eorum 5:10 miserunt ergo arcam Dei in Ac- 5:10 Therefore they sent the ark of God to Ekron. And it

came to pass, as the ark of God came to Ekron, that the Ekronites cried out, saying, They have brought about the

caron cumque uenisset arca Dei in Accaron exclamauerunt ark of the God of Israel to us, to slay us and our people. Accaronitae dicentes adduxerunt ad nos arcam Dei Israhel ut interficiat nos et populum nostrum 5:11 miserunt itaque 5:11 So they sent and gathered together all the lords of

the Philistines, and said, Send away the ark of the God of Israel, and let it go again to his own place, that it slay us not, and our people: for there was a deadly destruction throughout all the city; the hand of God was very heavy there.

et congregauerunt omnes satrapas Philisthinorum qui dixerunt dimittite arcam Dei Israhel et reuertatur in locum suum et non interficiat nos cum populo nostro 5:12 fiebat enim 5:12 And the men that died not were smitten with the

emerods: and the cry of the city went up to heaven.

pauor mortis in singulis urbibus et grauissima ualde manus Dei uiri quoque qui mortui non fuerant percutiebantur in secretiori parte natium et ascendebat ululatus uniuscuiusque ciuitatis in caelum

6:1 fuit ergo arca Domini in regione Philisthinorum se- 6:1 And the ark of the LORD was in the country of the

Philistines seven months.

ptem mensibus 6:2 et uocauerunt Philisthim sacerdotes et 6:2 And the Philistines called for the priests and the diviners, saying, What shall we do to the ark of the LORD? tell us wherewith we shall send it to his place.

diuinos dicentes quid faciemus de arca Dei indicate nobis quomodo remittemus eam in locum suum qui dixerunt 6:3 6:3 And they said, If ye send away the ark of the God

of Israel, send it not empty; but in any wise return him a trespass offering: then ye shall be healed, and it shall be

si remittitis arcam Dei Israhel nolite dimittere eam uacuam known to you why his hand is not removed from you. sed quod debetis reddite ei pro peccato et tunc curabimini et scietis quare non recedat manus eius a uobis 6:4 qui dixe- 6:4 Then said they, What shall be the trespass offering

which we shall return to him? They answered, Five golden emerods, and five golden mice, according to the number of the lords of the Philistines: for one plague was on you all, and on your lords.

runt quid est quod pro delicto reddere debeamus ei responderuntque illi 6:5 iuxta numerum prouinciarum Philisthim 6:5 Wherefore ye shall make images of your emerods, and

images of your mice that mar the land; and ye shall give glory unto the God of Israel: peradventure he will lighten his hand from off you, and from off your gods, and from off your land.

quinque anos aureos facietis et quinque mures aureos quia

Biblia Sacra Vulgata

plaga una fuit omnibus uobis et satrapis uestris facietisque similitudines anorum uestrorum et similitudines murium qui demoliti sunt terram et dabitis Deo Israhel gloriam si forte releuet manum suam a uobis et a diis uestris et a terra uestra

6:6 Wherefore then do ye harden your hearts, as the Egyp- 6:6 quare grauatis corda uestra sicut adgrauauit Aegyptus et

tians and Pharaoh hardened their hearts? when he had wrought wonderfully among them, did they not let the

people go, and they departed? Pharao cor suum nonne postquam percussus est tunc dimi-6:7 Now therefore make a new cart, and take two milch sit eos et abierunt 6:7 nunc ergo arripite et facite plaustrum

kine, on which there hath come no yoke, and tie the kine to the cart, and bring their calves home from them:

nouum unum et duas uaccas fetas quibus non est inpositum iugum iungite in plaustro et recludite uitulos earum domi

6:8 And take the ark of the LORD, and lay it upon the 6:8 tolletisque arcam Domini et ponetis in plaustro et uasa

cart; and put the jewels of gold, which ye return him for a trespass offering, in a coffer by the side thereof; and send

it away, that it may go. aurea quae exsoluistis ei pro delicto ponetis in capsella ad 6:9 And see, if it goeth up by the way of his own coast to latus eius et dimittite eam ut uadat 6:9 et aspicietis et si

Bethshemesh, then he hath done us this great evil: but if not, then we shall know that it is not his hand that smote

us: it was a chance that happened to us. quidem per uiam finium suorum ascenderit contra Bethsames ipse fecit nobis malum hoc grande sin autem minime sciemus quia nequaquam manus eius tetigit nos sed casu ac-

6:10 And the men did so; and took two milch kine, and cidit 6:10 fecerunt ergo illi hoc modo et tollentes duas uac-

tied them to the cart, and shut up their calves at home:

cas quae lactabant uitulos iunxerunt ad plaustrum uitulosque

6:11 And they laid the ark of the LORD upon the cart, and earum concluserunt domi 6:11 et posuerunt arcam Dei su-

the coffer with the mice of gold and the images of their emerods.

per plaustrum et capsellam quae habebat mures aureos et

6:12 And the kine took the straight way to the way of similitudinem anorum 6:12 ibant autem in directum uaccae

Bethshemesh, and went along the highway, lowing as they went, and turned not aside to the right hand or to the left; and the lords of the Philistines went after them unto the border of Bethshemesh.

per uiam quae ducit Bethsames et itinere uno gradiebantur pergentes et mugientes et non declinabant neque ad dextram neque ad sinistram sed et satrapae Philisthinorum sequeban-

6:13 And they of Bethshemesh were reaping their wheat tur usque ad terminos Bethsames 6:13 porro Bethsamitae

harvest in the valley: and they lifted up their eyes, and saw the ark, and rejoiced to see it.

metebant triticum in ualle et eleuantes oculos uiderunt ar-

6:14 And the cart came into the field of Joshua, a cam et gauisi sunt cum uidissent 6:14 et plaustrum uenit in

Bethshemite, and stood there, where there was a great stone: and they clave the wood of the cart, and offered

the kine a burnt offering unto the LORD. agrum Iosue Bethsamitae et stetit ibi erat autem ibi lapis ma-

gnus et conciderunt ligna plaustri uaccasque inposuerunt su-

6:15 And the Levites took down the ark of the LORD, per ea holocaustum Domino 6:15 Leuitae autem deposue-

and the coffer that was with it, wherein the jewels of gold were, and put them on the great stone: and the men of Bethshemesh offered burnt offerings and sacrificed sacrifices the same day unto the LORD.

runt arcam Dei et capsellam quae erat iuxta eam in qua erant uasa aurea et posuerunt super lapidem grandem uiri autem bethsamitae obtulerunt holocausta et immolauerunt uictimas

6:16 And when the five lords of the Philistines had seen it, in die illa Domino 6:16 et quinque satrapae Philisthinorum

they returned to Ekron the same day.

Liber I Samuelis

uiderunt et reuersi sunt in Accaron in die illa 6:17 hii sunt 6:17 And these are the golden emerods which the

Philistines returned for a trespass offering unto the LORD; for Ashdod one, for Gaza one, for Askelon one, for Gath

autem ani aurei quos reddiderunt Philisthim pro delicto Do- one, for Ekron one; mino Azotus unum Gaza unum Ascalon unum Geth unum

Accaron unum 6:18 et mures aureos secundum numerum 6:18 And the golden mice, according to the number of all the cities of the Philistines belonging to the five lords, both of fenced cities, and of country villages, even unto the great stone of Abel, whereon they set down the ark of the LORD: which stone remaineth unto this day in the field of Joshua, the Bethshemite.

urbium Philisthim quinque prouinciarum ab urbe murata usque ad uillam quae erat absque muro et usque ad Abel magnum super quem posuerunt arcam Domini quae erat usque in illa die in agro Iosue Bethsamitis 6:19 percussit autem 6:19 And he smote the men of Bethshemesh, because they

had looked into the ark of the LORD, even he smote of the people fifty thousand and threescore and ten men: and the people lamented, because the LORD had smitten many of the people with a great slaughter.

de uiris bethsamitibus eo quod uidissent arcam Domini et percussit de populo septuaginta uiros et quinquaginta milia plebis luxitque populus quod percussisset Dominus plebem plaga magna 6:20 et dixerunt uiri bethsamitae quis pote- 6:20 And the men of Bethshemesh said, Who is able to

stand before this holy LORD God? and to whom shall he go up from us?

rit stare in conspectu Domini Dei sancti huius et ad quem ascendet a nobis 6:21 miseruntque nuntios ad habitatores 6:21 And they sent messengers to the inhabitants of Kir-

jathjearim, saying, The Philistines have brought again the ark of the LORD; come ye down, and fetch it up to you.

Cariathiarim dicentes reduxerunt Philisthim arcam Domini descendite et ducite eam ad uos

7:1 uenerunt ergo uiri Cariathiarim et duxerunt arcam Do- 7:1 And the men of Kirjathjearim came, and fetched up

the ark of the LORD, and brought it into the house of Abinadab in the hill, and sanctified Eleazar his son to keep the

mini et intulerunt eam in domum Abinadab in Gabaa Elea- ark of the LORD. zarum autem filium eius sanctificauerunt ut custodiret arcam Domini 7:2 et factum est ex qua die mansit arca in Caria- 7:2 And it came to pass, while the ark abode in Kirjath-

jearim, that the time was long; for it was twenty years: and all the house of Israel lamented after the LORD.

thiarim multiplicati sunt dies erat quippe iam annus uicesimus et requieuit omnis domus Israhel post Dominum 7:3 7:3 And Samuel spake unto all the house of Israel, say-

ing, If ye do return unto the LORD with all your hearts, then put away the strange gods and Ashtaroth from among you, and prepare your hearts unto the LORD, and serve him only: and he will deliver you out of the hand of the Philistines.

ait autem Samuhel ad uniuersam domum Israhel dicens si in toto corde uestro reuertimini ad Dominum auferte deos alienos de medio uestrum et Astharoth et praeparate corda uestra Domino et seruite ei soli et eruet uos de manu Philisthim 7:4 abstulerunt ergo filii Israhel Baalim et Astharoth 7:4 Then the children of Israel did put away Baalim and

Ashtaroth, and served the LORD only.

et seruierunt Domino soli 7:5 dixit autem Samuhel congre- 7:5 And Samuel said, Gather all Israel to Mizpeh, and I

will pray for you unto the LORD.

gate uniuersum Israhel in Masphat ut orem pro uobis Domi-

num 7:6 et conuenerunt in Masphat hauseruntque aquam et 7:6 And they gathered together to Mizpeh, and drew water, and poured it out before the LORD, and fasted on that day, and said there, We have sinned against the LORD. And

effuderunt in conspectu Domini et ieiunauerunt in die illa et Samuel judged the children of Israel in Mizpeh. dixerunt ibi peccauimus Domino iudicauitque Samuhel filios Israhel in Masphat 7:7 et audierunt Philisthim quod con- 7:7 And when the Philistines heard that the children of

Israel were gathered together to Mizpeh, the lords of the Philistines went up against Israel. And when the children of Israel heard it, they were afraid of the Philistines.

Biblia Sacra Vulgata

gregati essent filii Israhel in Masphat et ascenderunt satrapae Philisthinorum ad Israhel quod cum audissent filii Israhel ti-

7:8 And the children of Israel said to Samuel, Cease not muerunt a facie Philisthinorum 7:8 dixeruntque ad Samu-

to cry unto the LORD our God for us, that he will save us out of the hand of the Philistines.

hel ne cesses pro nobis clamare ad Dominum Deum nostrum

7:9 And Samuel took a sucking lamb, and offered it for a ut saluet nos de manu Philisthinorum 7:9 tulit autem Samu-

burnt offering wholly unto the LORD: and Samuel cried unto the LORD for Israel; and the LORD heard him.

hel agnum lactantem unum et obtulit illum holocaustum integrum Domino et clamauit Samuhel ad Dominum pro Isra-

7:10 And as Samuel was offering up the burnt offering, hel et exaudiuit eum Dominus 7:10 factum est ergo cum Sa-

the Philistines drew near to battle against Israel: but the LORD thundered with a great thunder on that day upon the Philistines, and discomfited them; and they were smitten before Israel.

muhel offerret holocaustum Philistheos inire proelium contra Israhel intonuit autem Dominus fragore magno in die illa super Philisthim et exterruit eos et caesi sunt a filiis Israhel

7:11 And the men of Israel went out of Mizpeh, and pur- 7:11 egressique uiri Israhel de Masphat persecuti sunt Phi-

sued the Philistines, and smote them, until they came under Bethcar.

listheos et percusserunt eos usque ad locum qui erat subter

7:12 Then Samuel took a stone, and set it between Mizpeh Bethchar 7:12 tulit autem Samuhel lapidem unum et poand Shen, and called the name of it Ebenezer, saying, Hitherto hath the LORD helped us.

suit eum inter Masphat et inter Sen et uocauit nomen eius lapis Adiutorii dixitque hucusque auxiliatus est nobis Domi-

7:13 So the Philistines were subdued, and they came no nus 7:13 et humiliati sunt Philisthim nec adposuerunt ultra

more into the coast of Israel: and the hand of the LORD was against the Philistines all the days of Samuel.

ut uenirent in terminos Israhel facta est itaque manus Do-

7:14 And the cities which the Philistines had taken from mini super Philistheos cunctis diebus Samuhel 7:14 et red-

Israel were restored to Israel, from Ekron even unto Gath; and the coasts thereof did Israel deliver out of the hands of the Philistines. And there was peace between Israel and the Amorites.

ditae sunt urbes quas tulerant Philisthim ab Israhel Israheli ab Accaron usque Geth et terminos suos liberauit Israhel de manu Philisthinorum eratque pax inter Israhel et Amorreum

7:15 And Samuel judged Israel all the days of his life. 7:15 iudicabat quoque Samuhel Israhel cunctis diebus uitae 7:16 And he went from year to year in circuit to Bethel, suae 7:16 et ibat per singulos annos circumiens Bethel et

and Gilgal, and Mizpeh, and judged Israel in all those places.

Galgal et Masphat et iudicabat Israhelem in supradictis locis

7:17 And his return was to Ramah; for there was his house; 7:17 reuertebaturque in Ramatha ibi enim erat domus eius et

and there he judged Israel; and there he built an altar unto the LORD.

ibi iudicabat Israhelem aedificauit etiam ibi altare Domino

8:1 And it came to pass, when Samuel was old, that he 8:1 factum est autem cum senuisset Samuhel posuit filios

made his sons judges over Israel.

8:2 Now the name of his firstborn was Joel; and the name suos iudices Israhel 8:2 fuitque nomen filii eius primoge-

of his second, Abiah: they were judges in Beersheba.

8:3 And his sons walked not in his ways, but turned aside niti Iohel et nomen secundi Abia iudicum in Bersabee 8:3

after lucre, and took bribes, and perverted judgment.

et non ambulauerunt filii illius in uiis eius sed declinauerunt post auaritiam acceperuntque munera et peruerterunt iudi-

8:4 Then all the elders of Israel gathered themselves to- cium 8:4 congregati ergo uniuersi maiores natu Israhel ue-

gether, and came to Samuel unto Ramah,

Liber I Samuelis

nerunt ad Samuhel in Ramatha 8:5 dixeruntque ei ecce tu 8:5 And said unto him, Behold, thou art old, and thy sons

walk not in thy ways: now make us a king to judge us like all the nations.

senuisti et filii tui non ambulant in uiis tuis constitue nobis regem ut iudicet nos sicut uniuersae habent nationes 8:6 8:6 But the thing displeased Samuel, when they said, Give

us a king to judge us. And Samuel prayed unto the LORD.

displicuitque sermo in oculis Samuhelis eo quod dixissent da nobis regem ut iudicet nos et orauit Samuhel Dominum 8:7 dixit autem Dominus ad Samuhel audi uocem populi in 8:7 And the LORD said unto Samuel, Hearken unto the

voice of the people in all that they say unto thee: for they have not rejected thee, but they have rejected me, that I

omnibus quae loquuntur tibi non enim te abiecerunt sed me should not reign over them.

ne regnem super eos 8:8 iuxta omnia opera sua quae fece- 8:8 According to all the works which they have done since the day that I brought them up out of Egypt even unto this day, wherewith they have forsaken me, and served other

runt a die qua eduxi eos de Aegypto usque ad diem hanc gods, so do they also unto thee. sicut dereliquerunt me et seruierunt diis alienis sic faciunt

etiam tibi 8:9 nunc ergo audi uocem eorum uerumtamen 8:9 Now therefore hearken unto their voice: howbeit yet protest solemnly unto them, and shew them the manner of the king that shall reign over them.

contestare eos et praedic eis ius regis qui regnaturus est super eos 8:10 dixit itaque Samuhel omnia uerba Domini ad 8:10 And Samuel told all the words of the LORD unto the

people that asked of him a king.

populum qui petierat a se regem 8:11 et ait hoc erit ius re- 8:11 And he said, This will be the manner of the king that

shall reign over you: He will take your sons, and appoint them for himself, for his chariots, and to be his horsemen;

gis qui imperaturus est uobis filios uestros tollet et ponet in and some shall run before his chariots. curribus suis facietque sibi equites et praecursores quadrigarum suarum 8:12 et constituet sibi tribunos et centuriones et 8:12 And he will appoint him captains over thousands, and

captains over fifties; and will set them to ear his ground, and to reap his harvest, and to make his instruments of

aratores agrorum suorum et messores segetum et fabros ar- war, and instruments of his chariots. morum et curruum suorum 8:13 filias quoque uestras faciet 8:13 And he will take your daughters to be confectionar-

ies, and to be cooks, and to be bakers.

sibi unguentarias et focarias et panificas 8:14 agros quoque 8:14 And he will take your fields, and your vineyards, and

your oliveyards, even the best of them, and give them to his servants.

uestros et uineas et oliueta optima tollet et dabit seruis suis 8:15 sed et segetes uestras et uinearum rediius addecimabit 8:15 And he will take the tenth of your seed, and of your

vineyards, and give to his officers, and to his servants.

ut det eunuchis et famulis suis 8:16 seruos etiam uestros 8:16 And he will take your menservants, and your maid-

servants, and your goodliest young men, and your asses, and put them to his work.

et ancillas et iuuenes optimos et asinos auferet et ponet in opere suo 8:17 greges uestros addecimabit uosque eritis ei 8:17 He will take the tenth of your sheep: and ye shall be

his servants.

serui 8:18 et clamabitis in die illa a facie regis uestri quem 8:18 And ye shall cry out in that day because of your king

which ye shall have chosen you; and the LORD will not hear you in that day.

elegistis uobis et non exaudiet uos Dominus in die illa 8:19

8:19 Nevertheless the people refused to obey the voice of Samuel; and they said, Nay; but we will have a king over us;

noluit autem populus audire uocem Samuhel sed dixerunt nequaquam rex enim erit super nos 8:20 et erimus nos quo- 8:20 That we also may be like all the nations; and that our

king may judge us, and go out before us, and fight our battles.

que sicut omnes gentes et iudicabit nos rex noster et egredietur ante nos et pugnabit bella nostra pro nobis 8:21 et 8:21 And Samuel heard all the words of the people, and

he rehearsed them in the ears of the LORD.

audiuit Samuhel omnia uerba populi et locutus est ea in auribus Domini 8:22 dixit autem Dominus ad Samuhel audi 8:22 And the LORD said to Samuel, Hearken unto their

voice, and make them a king. And Samuel said unto the men of Israel, Go ye every man unto his city.

Biblia Sacra Vulgata

uocem eorum et constitue super eos regem et ait Samuhel ad uiros Israhel uadat unusquisque in ciuitatem suam

9:1 Now there was a man of Benjamin, whose name was 9:1 et erat uir de Beniamin nomine Cis filius Abihel filii

Kish, the son of Abiel, the son of Zeror, the son of Bechorath, the son of Aphiah, a Benjamite, a mighty man of

power. Seror filii Bechoreth filii Afia filii uiri Iemini fortis robore 9:2 And he had a son, whose name was Saul, a choice 9:2 et erat ei filius uocabulo Saul electus et bonus et non erat

young man, and a goodly: and there was not among the children of Israel a goodlier person than he: from his shoulders and upward he was higher than any of the people.

uir de filiis Israhel melior illo ab umero et sursum eminebat

9:3 And the asses of Kish Saul’s father were lost. And super omnem populum 9:3 perierant autem asinae Cis pat-

Kish said to Saul his son, Take now one of the servants with thee, and arise, go seek the asses.

ris Saul et dixit Cis ad Saul filium suum tolle tecum unum de pueris et consurgens uade et quaere asinas qui cum transis-

9:4 And he passed through mount Ephraim, and passed sent per montem Ephraim 9:4 et per terram Salisa et non in-

through the land of Shalisha, but they found them not: then they passed through the land of Shalim, and there they were not: and he passed through the land of the Benjamites, but they found them not.

uenissent transierunt etiam per terram Salim et non erant sed

9:5 And when they were come to the land of Zuph, Saul et per terram Iemini et minime reppererunt 9:5 cum autem

said to his servant that was with him, Come, and let us return; lest my father leave caring for the asses, and take

thought for us. uenissent in terram Suph dixit Saul ad puerum suum qui erat

cum eo ueni et reuertamur ne forte dimiserit pater meus asi-

9:6 And he said unto him, Behold now, there is in this city nas et sollicitus sit pro nobis 9:6 qui ait ei ecce est uir Dei

a man of God, and he is an honourable man; all that he saith cometh surely to pass: now let us go thither; perad-

venture he can shew us our way that we should go. in ciuitate hac uir nobilis omne quod loquitur absque am-

biguitate uenit nunc ergo eamus illuc si forte indicet nobis

9:7 Then said Saul to his servant, But, behold, if we go, de uia nostra propter quam uenimus 9:7 dixitque Saul ad what shall we bring the man? for the bread is spent in our vessels, and there is not a present to bring to the man of

God: what have we? puerum suum ecce ibimus quid feremus ad uirum panis de-

fecit in sitarciis nostris et sportulam non habemus ut demus

9:8 And the servant answered Saul again, and said, Be- homini Dei nec quicquam aliud 9:8 rursum puer respondit

hold, I have here at hand the fourth part of a shekel of silver: that will I give to the man of God, to tell us our

way. Sauli et ait ecce inuenta est in manu mea quarta pars state-

ris argenti demus homini Dei ut indicet nobis uiam nostram

9:9 (Beforetime in Israel, when a man went to enquire of 9:9 olim in Israhel sic loquebatur unusquisque uadens consu-

God, thus he spake, Come, and let us go to the seer: for he that is now called a Prophet was beforetime called a Seer.)

lere Deum uenite et eamus ad uidentem qui enim propheta

9:10 Then said Saul to his servant, Well said; come, let us dicitur hodie uocabatur olim uidens 9:10 et dixit Saul ad go. So they went unto the city where the man of God was.

puerum suum optimus sermo tuus ueni eamus et ierunt in

9:11 And as they went up the hill to the city, they found ciuitatem in qua erat uir Dei 9:11 cumque ascenderent cli-

young maidens going out to draw water, and said unto them, Is the seer here?

uum ciuitatis inuenerunt puellas egredientes ad hauriendam

9:12 And they answered them, and said, He is; behold, he aquam et dixerunt eis num hic est uidens 9:12 quae respon-

is before you: make haste now, for he came to day to the city; for there is a sacrifice of the people to day in the high

place: dentes dixerunt illis hic est ecce ante te festina nunc hodie

enim uenit in ciuitate quia sacrificium est hodie populo in

9:13 As soon as ye be come into the city, ye shall straight- excelso 9:13 ingredientes urbem statim inuenietis eum an-

way find him, before he go up to the high place to eat: for the people will not eat until he come, because he doth bless the sacrifice; and afterwards they eat that be bidden. Now therefore get you up; for about this time ye shall find him.

CDLXIV

Liber I Samuelis

tequam ascendat excelsum ad uescendum neque enim comesurus est populus donec ille ueniat quia ipse benedicit hostiae et deinceps comedunt qui uocati sunt nunc ergo conscendite quia hodie repperietis eum 9:14 et ascenderunt in 9:14 And they went up into the city: and when they were

come into the city, behold, Samuel came out against them, for to go up to the high place.

ciuitatem cumque illi ambularent in medio urbis apparuit Samuhel egrediens obuiam eis ut ascenderet in excelsum 9:15 9:15 Now the LORD had told Samuel in his ear a day be-

fore Saul came, saying,

Dominus autem reuelauerat auriculam Samuhel ante unam diem quam ueniret Saul dicens 9:16 hac ipsa quae nunc est 9:16 To morrow about this time I will send thee a man out

of the land of Benjamin, and thou shalt anoint him to be captain over my people Israel, that he may save my people out of the hand of the Philistines: for I have looked upon my people, because their cry is come unto me.

hora cras mittam ad te uirum de terra Beniamin et ungues eum ducem super populum meum Israhel et saluabit populum meum de manu Philisthinorum quia respexi populum meum uenit enim clamor eorum ad me 9:17 cumque as- 9:17 And when Samuel saw Saul, the LORD said unto

him, Behold the man whom I spake to thee of! this same shall reign over my people.

pexisset Samuhel Saulem Dominus ait ei ecce uir quem dixeram tibi iste dominabitur populo meo 9:18 accessit autem 9:18 Then Saul drew near to Samuel in the gate, and said,

Tell me, I pray thee, where the seer’s house is.

Saul ad Samuhelem in medio portae et ait indica oro mihi ubi est domus uidentis 9:19 et respondit Samuhel Sauli di- 9:19 And Samuel answered Saul, and said, I am the seer:

go up before me unto the high place; for ye shall eat with me to day, and to morrow I will let thee go, and will tell

cens ego sum uidens ascende ante me in excelsum ut come- thee all that is in thine heart. datis mecum hodie et dimittam te mane et omnia quae sunt in corde tuo indicabo tibi 9:20 et de asinis quas perdidisti 9:20 And as for thine asses that were lost three days ago,

set not thy mind on them; for they are found. And on whom is all the desire of Israel? Is it not on thee, and on

nudius tertius ne sollicitus sis quia inuentae sunt et cuius all thy father’s house? erunt optima quaeque Israhel nonne tibi et omni domui patris tui 9:21 respondens autem Saul ait numquid non filius 9:21 And Saul answered and said, Am not I a Benjamite,

of the smallest of the tribes of Israel? and my family the least of all the families of the tribe of Benjamin? where-

Iemini ego sum de minima tribu Israhel et cognatio mea no- fore then speakest thou so to me? uissima inter omnes familias de tribu Beniamin quare ergo locutus es mihi sermonem istum 9:22 adsumens itaque Sa- 9:22 And Samuel took Saul and his servant, and brought

them into the parlour, and made them sit in the chiefest place among them that were bidden, which were about

muhel Saulem et puerum eius introduxit eos in triclinium thirty persons. et dedit eis locum in capite eorum qui fuerant inuitati erant enim quasi triginta uiri 9:23 dixitque Samuhel coco da par- 9:23 And Samuel said unto the cook, Bring the portion

which I gave thee, of which I said unto thee, Set it by thee.

tem quam dedi tibi et praecepi ut reponeres seorsum apud te 9:24 leuauit autem cocus armum et posuit ante Saul dixitque 9:24 And the cook took up the shoulder, and that which

was upon it, and set it before Saul. And Samuel said, Behold that which is left! set it before thee, and eat: for unto this time hath it been kept for thee since I said, I have invited the people. So Saul did eat with Samuel that day.

Samuhel ecce quod remansit pone ante te et comede quia de industria seruatum est tibi quando populum uocaui et comedit Saul cum Samuhel in die illa 9:25 et descenderunt de 9:25 And when they were come down from the high place

into the city, Samuel communed with Saul upon the top of the house.

Biblia Sacra Vulgata

9:26 And they arose early: and it came to pass about the excelso in oppidum et locutus est cum Saul in solario 9:26

spring of the day, that Samuel called Saul to the top of the house, saying, Up, that I may send thee away. And Saul arose, and they went out both of them, he and Samuel, abroad.

cumque mane surrexissent et iam dilucesceret uocauit Samuhel Saul in solarium dicens surge ut dimittam te et surrexit Saul egressique sunt ambo ipse uidelicet et Samuhel

9:27 And as they were going down to the end of the city, 9:27 cumque descenderent in extrema parte ciuitatis Samu-

Samuel said to Saul, Bid the servant pass on before us, (and he passed on), but stand thou still a while, that I may

shew thee the word of God. hel dixit ad Saul dic puero ut antecedat nos et transeat tu autem subsiste paulisper ut indicem tibi uerbum Domini

10:1 Then Samuel took a vial of oil, and poured it upon his 10:1 tulit autem Samuhel lenticulam olei et effudit super

head, and kissed him, and said, Is it not because the LORD hath anointed thee to be captain over his inheritance?

caput eius et deosculatus eum ait ecce unxit te Dominus su-

10:2 When thou art departed from me to day, then thou per hereditatem suam in principem 10:2 cum abieris hodie

shalt find two men by Rachel’s sepulchre in the border of Benjamin at Zelzah; and they will say unto thee, The asses which thou wentest to seek are found: and, lo, thy father hath left the care of the asses, and sorroweth for you, saying, What shall I do for my son?

a me inuenies duos uiros iuxta sepulchrum Rachel in finibus Beniamin in meridie dicentque tibi inuentae sunt asinae ad quas ieras perquirendas et intermissis pater tuus asinis sol-

10:3 Then shalt thou go on forward from thence, and thou licitus est pro uobis et dicit quid faciam de filio meo 10:3

shalt come to the plain of Tabor, and there shall meet thee three men going up to God to Bethel, one carrying three kids, and another carrying three loaves of bread, and another carrying a bottle of wine:

cumque abieris inde et ultra transieris et ueneris ad quercum Thabor inuenient te ibi tres uiri ascendentes ad Deum in Bethel unus portans tres hedos et alius tres tortas panis

10:4 And they will salute thee, and give thee two loaves of et alius portans lagoenam uini 10:4 cumque te salutauerint

bread; which thou shalt receive of their hands.

10:5 After that thou shalt come to the hill of God, where dabunt tibi duos panes et accipies de manu eorum 10:5 post

is the garrison of the Philistines: and it shall come to pass, when thou art come thither to the city, that thou shalt meet a company of prophets coming down from the high place with a psaltery, and a tabret, and a pipe, and a harp, before them; and they shall prophesy:

haec uenies in collem Domini ubi est statio Philisthinorum et cum ingressus fueris ibi urbem obuiam habebis gregem prophetarum descendentium de excelso et ante eos psalterium et tympanum et tibiam et citharam ipsosque prophe-

10:6 And the Spirit of the LORD will come upon thee, tantes 10:6 et insiliet in te spiritus Domini et prophetabis

and thou shalt prophesy with them, and shalt be turned into another man.

cum eis et mutaberis in uirum alium 10:7 quando ergo eue-

10:7 And let it be, when these signs are come unto thee, that thou do as occasion serve thee; for God is with thee.

nerint signa haec omnia tibi fac quaecumque inuenerit ma-

10:8 And thou shalt go down before me to Gilgal; and, be- nus tua quia Dominus tecum est 10:8 et descendes ante me hold, I will come down unto thee, to offer burnt offerings, and to sacrifice sacrifices of peace offerings: seven days shalt thou tarry, till I come to thee, and shew thee what thou shalt do.

in Galgala ego quippe descendam ad te ut offeras oblationem et immoles uictimas pacificas septem diebus expecta-

10:9 And it was so, that when he had turned his back to go bis donec ueniam ad te et ostendam tibi quae facias 10:9

from Samuel, God gave him another heart: and all those signs came to pass that day.

itaque cum auertisset umerum suum ut abiret a Samuhele inmutauit ei Deus cor aliud et uenerunt omnia signa haec

10:10 And when they came thither to the hill, behold, a in die illa 10:10 ueneruntque ad praedictum collem et ecce

company of prophets met him; and the Spirit of God came upon him, and he prophesied among them.

Liber I Samuelis

cuneus prophetarum obuius ei et insiliuit super eum spiritus Dei et prophetauit in medio eorum 10:11 uidentes autem 10:11 And it came to pass, when all that knew him before-

time saw that, behold, he prophesied among the prophets, then the people said one to another, What is this that is come unto the son of Kish? Is Saul also among the prophets?

omnes qui nouerant eum heri et nudius tertius quod esset cum prophetis et prophetaret dixerunt ad inuicem quaenam res accidit filio Cis num et Saul in prophetis 10:12 respon- 10:12 And one of the same place answered and said, But

who is their father? Therefore it became a proverb, Is Saul also among the prophets?

ditque alius ad alterum dicens et quis pater eorum propterea uersum est in prouerbium num et Saul inter prophetas 10:13 cessauit autem prophetare et uenit ad excelsum 10:14 10:13 And when he had made an end of prophesying, he

came to the high place.

10:14 And Saul’s uncle said unto him and to his servant, Whither went ye? And he said, To seek the asses: and when we saw that they were no where, we came to Samuel.

dixitque patruus Saul ad eum et ad puerum eius quo abistis qui responderunt quaerere asinas quas cum non repperissemus uenimus ad Samuhelem 10:15 et dixit ei patruus suus 10:15 And Saul’s uncle said, Tell me, I pray thee, what

Samuel said unto you.

indica mihi quid dixerit tibi Samuhel 10:16 et ait Saul ad 10:16 And Saul said unto his uncle, He told us plainly that

the asses were found. But of the matter of the kingdom, whereof Samuel spake, he told him not.

patruum suum indicauit nobis quia inuentae essent asinae de sermone autem regni non indicauit ei quem locutus illi fuerat Samuhel 10:17 et conuocauit Samuhel populum ad 10:17 And Samuel called the people together unto the

LORD to Mizpeh;

Dominum in Maspha 10:18 et ait ad filios Israhel haec di- 10:18 And said unto the children of Israel, Thus saith the LORD God of Israel, I brought up Israel out of Egypt, and delivered you out of the hand of the Egyptians, and out of

cit Dominus Deus Israhel ego eduxi Israhel de Aegypto et the hand of all kingdoms, and of them that oppressed you: erui uos de manu Aegyptiorum et de manu omnium regum qui adfligebant uos 10:19 uos autem hodie proiecistis Deum 10:19 And ye have this day rejected your God, who him-

self saved you out of all your adversities and your tribulations; and ye have said unto him, Nay, but set a king over us. Now therefore present yourselves before the LORD by your tribes, and by your thousands.

uestrum qui solus saluauit uos de uniuersis malis et tribulationibus uestris et dixistis nequaquam sed regem constitue super nos nunc ergo state coram Domino per tribus uestras et per familias 10:20 et adplicuit Samuhel omnes tribus Is- 10:20 And when Samuel had caused all the tribes of Israel

to come near, the tribe of Benjamin was taken.

rahel et cecidit sors tribus Beniamin 10:21 et adplicuit tri- 10:21 When he had caused the tribe of Benjamin to come

near by their families, the family of Matri was taken, and Saul the son of Kish was taken: and when they sought him,

bum Beniamin et cognationes eius et cecidit cognatio Metri he could not be found. et peruenit usque ad Saul filium Cis quaesierunt ergo eum et non est inuentus 10:22 et consuluerunt post haec Domi- 10:22 Therefore they enquired of the LORD further, if the

man should yet come thither. And the LORD answered, Behold he hath hid himself among the stuff.

num utrumnam uenturus esset illuc responditque Dominus ecce absconditus est domi 10:23 cucurrerunt itaque et tule- 10:23 And they ran and fetched him thence: and when he

stood among the people, he was higher than any of the people from his shoulders and upward.

runt eum inde stetitque in medio populi et altior fuit uniuerso populo ab umero et sursum 10:24 et ait Samuhel ad omnem 10:24 And Samuel said to all the people, See ye him whom

the LORD hath chosen, that there is none like him among all the people? And all the people shouted, and said, God

populum certe uidetis quem elegit Dominus quoniam non sit save the king. similis ei in omni populo et clamauit cunctus populus et ait

Biblia Sacra Vulgata

10:25 Then Samuel told the people the manner of the king- uiuat rex 10:25 locutus est autem Samuhel ad populum ledom, and wrote it in a book, and laid it up before the LORD. And Samuel sent all the people away, every man

to his house. gem regni et scripsit in libro et reposuit coram Domino et

dimisit Samuhel omnem populum singulos in domum suam

10:26 And Saul also went home to Gibeah; and there went 10:26 sed et Saul abiit in domum suam in Gabaath et abiit

with him a band of men, whose hearts God had touched. 10:27 But the children of Belial said, How shall this man cum eo pars exercitus quorum tetigerat Deus corda 10:27 save us? And they despised him, and brought no presents. But he held his peace.

filii uero Belial dixerunt num saluare nos poterit iste et despexerunt eum et non adtulerunt ei munera ille uero dissimulabat se audire

11:1 Then Nahash the Ammonite came up, and encamped 11:1 ascendit autem Naas Ammonites et pugnare coepit

against Jabeshgilead: and all the men of Jabesh said unto Nahash, Make a covenant with us, and we will serve thee.

aduersus Iabesgalaad dixeruntque omnes uiri Iabes ad Naas

11:2 And Nahash the Ammonite answered them, On this habeto nos foederatos et seruiemus tibi 11:2 et respondit

condition will I make a covenant with you, that I may thrust out all your right eyes, and lay it for a reproach upon

all Israel. ad eos Naas Ammonites in hoc feriam uobiscum foedus ut

eruam omnium uestrum oculos dextros ponamque uos ob-

11:3 And the elders of Jabesh said unto him, Give us seven probrium in uniuerso Israhel 11:3 et dixerunt ad eum sedays’ respite, that we may send messengers unto all the coasts of Israel: and then, if there be no man to save us,

we will come out to thee. niores Iabes concede nobis septem dies ut mittamus nuntios

in uniuersos terminos Israhel et si non fuerit qui defendat

11:4 Then came the messengers to Gibeah of Saul, and nos egrediemur ad te 11:4 uenerunt ergo nuntii in Gabaath

told the tidings in the ears of the people: and all the people lifted up their voices, and wept.

Saulis et locuti sunt uerba audiente populo et leuauit omnis

11:5 And, behold, Saul came after the herd out of the field; populus uocem suam et fleuit 11:5 et ecce Saul ueniebat seand Saul said, What aileth the people that they weep? And they told him the tidings of the men of Jabesh.

quens boues de agro et ait quid habet populus quod plorat et

11:6 And the Spirit of God came upon Saul when he heard narrauerunt ei uerba uirorum Iabes 11:6 et insiliuit spiritus

those tidings, and his anger was kindled greatly.

Domini in Saul cum audisset uerba haec et iratus est furor

11:7 And he took a yoke of oxen, and hewed them in eius nimis 11:7 et adsumens utrumque bouem concidit in

pieces, and sent them throughout all the coasts of Israel by the hands of messengers, saying, Whosoever cometh not forth after Saul and after Samuel, so shall it be done unto his oxen. And the fear of the LORD fell on the people, and they came out with one consent.

frusta misitque in omnes terminos Israhel per manum nuntiorum dicens quicumque non exierit secutusque fuerit Saul et Samuhelem sic fiet bubus eius inuasit ergo timor Domini

11:8 And when he numbered them in Bezek, the children populum et egressi sunt quasi uir unus 11:8 et recensuit eos

of Israel were three hundred thousand, and the men of Judah thirty thousand.

in Bezec fueruntque filiorum Israhel trecenta milia uirorum

11:9 And they said unto the messengers that came, Thus autem Iuda triginta milia 11:9 et dixerunt nuntiis qui uene-

shall ye say unto the men of Jabeshgilead, To morrow, by that time the sun be hot, ye shall have help. And the messengers came and shewed it to the men of Jabesh; and they were glad.

rant sic dicetis uiris qui sunt in Iabesgalaad cras erit uobis salus cum incaluerit sol uenerunt ergo nuntii et adnuntia-

11:10 Therefore the men of Jabesh said, To morrow we uerunt uiris Iabes qui laetati sunt 11:10 et dixerunt mane

will come out unto you, and ye shall do with us all that seemeth good unto you.

exibimus ad uos et facietis nobis omne quod placuerit uobis

Liber I Samuelis

11:11 et factum est cum uenisset dies crastinus constituit Saul 11:11 And it was so on the morrow, that Saul put the peo-

ple in three companies; and they came into the midst of the host in the morning watch, and slew the Ammonites until the heat of the day: and it came to pass, that they which remained were scattered, so that two of them were not left together.

populum in tres partes et ingressus est media castra in uigilia matutina et percussit Ammon usque dum incalesceret dies reliqui autem dispersi sunt ita ut non relinquerentur in eis duo pariter 11:12 et ait populus ad Samuhel quis est iste 11:12 And the people said unto Samuel, Who is he that

said, Shall Saul reign over us? bring the men, that we may put them to death.

qui dixit Saul non regnabit super nos date uiros et interficiemus eos 11:13 et ait Saul non occidetur quisquam in die 11:13 And Saul said, There shall not a man be put to death

this day: for to day the LORD hath wrought salvation in Israel.

hac quia hodie fecit Dominus salutem in Israhel 11:14 dixit

11:14 Then said Samuel to the people, Come, and let us go to Gilgal, and renew the kingdom there.

autem Samuhel ad populum uenite et eamus in Galgala et innouemus ibi regnum 11:15 et perrexit omnis populus in 11:15 And all the people went to Gilgal; and there they

made Saul king before the LORD in Gilgal; and there they sacrificed sacrifices of peace offerings before the LORD;

Galgala et fecerunt ibi regem Saul coram Domino in Gal- and there Saul and all the men of Israel rejoiced greatly.

gala et immolauerunt ibi uictimas pacificas coram Domino et laetatus est ibi Saul et cuncti uiri Israhel nimis

12:1 dixit autem Samuhel ad uniuersum Israhel ecce au- 12:1 And Samuel said unto all Israel, Behold, I have hearkened unto your voice in all that ye said unto me, and have made a king over you.

diui uocem uestram iuxta omnia quae locuti estis ad me et constitui super uos regem 12:2 et nunc rex graditur ante uos 12:2 And now, behold, the king walketh before you: and

I am old and grayheaded; and, behold, my sons are with you: and I have walked before you from my childhood

ego autem senui et incanui porro filii mei uobiscum sunt ita- unto this day. que conuersatus coram uobis ab adulescentia mea usque ad diem hanc ecce praesto sum 12:3 loquimini de me coram 12:3 Behold, here I am: witness against me before the

LORD, and before his anointed: whose ox have I taken? or whose ass have I taken? or whom have I defrauded? whom have I oppressed? or of whose hand have I received any bribe to blind mine eyes therewith? and I will restore it you.

Domino et coram christo eius utrum bouem cuiusquam tulerim an asinum si quempiam calumniatus sum si oppressi aliquem si de manu cuiusquam munus accepi et contemnam illud hodie restituamque uobis 12:4 et dixerunt non es ca- 12:4 And they said, Thou hast not defrauded us, nor op-

pressed us, neither hast thou taken ought of any man’s hand.

lumniatus nos neque oppressisti neque tulisti de manu alicuius quippiam 12:5 dixitque ad eos testis Dominus aduer- 12:5 And he said unto them, The LORD is witness against

you, and his anointed is witness this day, that ye have not found ought in my hand. And they answered, He is wit-

sus uos et testis christus eius in die hac quia non inueneri- ness. tis in manu mea quippiam et dixerunt testis 12:6 et ait Sa- 12:6 And Samuel said unto the people, It is the LORD that

advanced Moses and Aaron, and that brought your fathers up out of the land of Egypt.

muhel ad populum Dominus qui fecit Mosen et Aaron et eduxit patres nostros de terra Aegypti 12:7 nunc ergo state 12:7 Now therefore stand still, that I may reason with you

before the LORD of all the righteous acts of the LORD, which he did to you and to your fathers.

ut iudicio contendam aduersum uos coram Domino de omnibus misericordiis Domini quas fecit uobiscum et cum patribus uestris 12:8 quomodo ingressus est Iacob in Aegyp- 12:8 When Jacob was come into Egypt, and your fathers

cried unto the LORD, then the LORD sent Moses and Aaron, which brought forth your fathers out of Egypt, and

tum et clamauerunt patres uestri ad Dominum et misit Do- made them dwell in this place.

Biblia Sacra Vulgata

minus Mosen et Aaron et eduxit patres uestros ex Aegypto

12:9 And when they forgat the LORD their God, he sold et conlocauit eos in loco hoc 12:9 qui obliti sunt Domini

them into the hand of Sisera, captain of the host of Hazor, and into the hand of the Philistines, and into the hand of

the king of Moab, and they fought against them. Dei sui et tradidit eos in manu Sisarae magistri militiae Asor

et in manu Philisthinorum et in manu regis Moab et pug-

12:10 And they cried unto the LORD, and said, We have nauerunt aduersum eos 12:10 postea autem clamauerunt ad

sinned, because we have forsaken the LORD, and have served Baalim and Ashtaroth: but now deliver us out of

the hand of our enemies, and we will serve thee. Dominum et dixerunt peccauimus quia dereliquimus Domi-

num et seruiuimus Baalim et Astharoth nunc ergo erue nos

12:11 And the LORD sent Jerubbaal, and Bedan, and de manu inimicorum nostrorum et seruiemus tibi 12:11 et

Jephthah, and Samuel, and delivered you out of the hand of your enemies on every side, and ye dwelled safe.

misit Dominus Hierobaal et Bedan et Ieptha et Samuhel et eruit uos de manu inimicorum uestrorum per circuitum et

12:12 And when ye saw that Nahash the king of the chil- habitastis confidenter 12:12 uidentes autem quod Naas rex

dren of Ammon came against you, ye said unto me, Nay; but a king shall reign over us: when the LORD your God

was your king. filiorum Ammon uenisset aduersum uos dixistis mihi nequa-

quam sed rex imperabit nobis cum Dominus Deus uester re-

12:13 Now therefore behold the king whom ye have cho- gnaret in uobis 12:13 nunc ergo praesto est rex uester quem

sen, and whom ye have desired! and, behold, the LORD hath set a king over you.

elegistis et petistis ecce dedit uobis Dominus regem 12:14

12:14 If ye will fear the LORD, and serve him, and obey his voice, and not rebel against the commandment of the LORD, then shall both ye and also the king that reigneth over you continue following the LORD your God:

si timueritis Dominum et seruieritis ei et audieritis uocem eius et non exasperaueritis os Domini eritis et uos et rex qui

12:15 But if ye will not obey the voice of the LORD, but imperat uobis sequentes Dominum Deum uestrum 12:15 si

rebel against the commandment of the LORD, then shall the hand of the LORD be against you, as it was against

your fathers. autem non audieritis uocem Domini sed exasperaueritis ser-

monem Domini erit manus Domini super uos et super patres

12:16 Now therefore stand and see this great thing, which uestros 12:16 sed et nunc state et uidete rem istam granthe LORD will do before your eyes.

12:17 Is it not wheat harvest to day? I will call unto the dem quam facturus est Dominus in conspectu uestro 12:17

LORD, and he shall send thunder and rain; that ye may perceive and see that your wickedness is great, which ye

have done in the sight of the LORD, in asking you a king. numquid non messis tritici est hodie inuocabo Dominum et dabit uoces et pluuias et scietis et uidebitis quia grande malum feceritis uobis in conspectu Domini petentes super uos

12:18 So Samuel called unto the LORD; and the LORD regem 12:18 et clamauit Samuhel ad Dominum et dedit Do-

sent thunder and rain that day: and all the people greatly feared the LORD and Samuel.

minus uoces et pluuiam in die illa 12:19 et timuit omnis

12:19 And all the people said unto Samuel, Pray for thy servants unto the LORD thy God, that we die not: for we have added unto all our sins this evil, to ask us a king.

populus nimis Dominum et Samuhelem dixitque uniuersus populus ad Samuhel ora pro seruis tuis ad Dominum Deum tuum ut non moriamur addidimus enim uniuersis peccatis

12:20 And Samuel said unto the people, Fear not: ye have nostris malum ut peteremus nobis regem 12:20 dixit autem

done all this wickedness: yet turn not aside from following the LORD, but serve the LORD with all your heart;

Samuhel ad populum nolite timere uos fecistis uniuersum malum hoc uerumtamen nolite recedere a tergo Domini et

Liber I Samuelis

seruite Domino in omni corde uestro 12:21 et nolite decli- 12:21 And turn ye not aside: for then should ye go after

vain things, which cannot profit nor deliver; for they are vain.

nare post uana quae non prodeeunt uobis neque eruent uos quia uana sunt 12:22 et non derelinquet Dominus populum 12:22 For the LORD will not forsake his people for his

great name’s sake: because it hath pleased the LORD to make you his people.

suum propter nomen suum magnum quia iurauit Dominus facere uos sibi populum 12:23 absit autem a me hoc pec- 12:23 Moreover as for me, God forbid that I should sin

against the LORD in ceasing to pray for you: but I will teach you the good and the right way:

catum in Domino ut cessem orare pro uobis et docebo uos uiam bonam et rectam 12:24 igitur timete Dominum et ser- 12:24 Only fear the LORD, and serve him in truth with all

your heart: for consider how great things he hath done for you.

uite ei in ueritate et ex toto corde uestro uidistis enim magnifica quae in uobis gesserit 12:25 quod si perseueraueritis 12:25 But if ye shall still do wickedly, ye shall be con-

sumed, both ye and your king.

in malitia et uos et rex uester pariter peribitis

13:1 filius unius anni Saul cum regnare coepisset duobus 13:1 Saul reigned one year; and when he had reigned two

years over Israel,

autem annis regnauit super Israhel 13:2 et elegit sibi Saul 13:2 Saul chose him three thousand men of Israel; whereof

two thousand were with Saul in Michmash and in mount Bethel, and a thousand were with Jonathan in Gibeah of Benjamin: and the rest of the people he sent every man to his tent.

tria milia de Israhel et erant cum Saul duo milia in Machmas et in monte Bethel mille autem cum Ionathan in Gabaath Beniamin porro ceterum populum remisit unumquemque in tabernacula sua 13:3 et percussit Ionathan stationem 13:3 And Jonathan smote the garrison of the Philistines

that was in Geba, and the Philistines heard of it. And Saul blew the trumpet throughout all the land, saying, Let the

Philisthim quae erat in Gabaa quod cum audissent Philist- Hebrews hear. him Saul cecinit bucina in omni terra dicens audiant Hebraei 13:4 et uniuersus Israhel audiuit huiuscemodi famam 13:4 And all Israel heard say that Saul had smitten a garri-

son of the Philistines, and that Israel also was had in abomination with the Philistines. And the people were called

percussit Saul stationem Philisthinorum et erexit se Israhel together after Saul to Gilgal. aduersum Philisthim clamauit ergo populus post Saul in Galgala 13:5 et Philisthim congregati sunt ad proeliandum con- 13:5 And the Philistines gathered themselves together to

fight with Israel, thirty thousand chariots, and six thousand horsemen, and people as the sand which is on the sea shore in multitude: and they came up, and pitched in Michmash, eastward from Bethaven.

tra Israhel triginta milia curruum et sex milia equitum et reliquum uulgus sicut harena quae est in litore maris plurima et ascendentes castrametati sunt in Machmas ad orientem Bethauen 13:6 quod cum uidissent uiri Israhel se in arto 13:6 When the men of Israel saw that they were in a strait,

(for the people were distressed,) then the people did hide themselves in caves, and in thickets, and in rocks, and in

sitos adflictus est enim populus absconderunt se in spelun- high places, and in pits. cis et in abditis in petris quoque et in antris et in cisternis 13:7 Hebraei autem transierunt Iordanem terram Gad et Ga- 13:7 And some of the Hebrews went over Jordan to the

land of Gad and Gilead. As for Saul, he was yet in Gilgal, and all the people followed him trembling.

laad cumque adhuc esset Saul in Galgal uniuersus populus perterritus est qui sequebatur eum 13:8 et expectauit sep- 13:8 And he tarried seven days, according to the set time

that Samuel had appointed: but Samuel came not to Gilgal; and the people were scattered from him.

tem diebus iuxta placitum Samuhel et non uenit Samuhel in Galgala dilapsusque est populus ab eo 13:9 ait ergo Saul 13:9 And Saul said, Bring hither a burnt offering to me,

and peace offerings. And he offered the burnt offering.

Biblia Sacra Vulgata

adferte mihi holocaustum et pacifica et obtulit holocaustum

13:10 And it came to pass, that as soon as he had made an 13:10 cumque conplesset offerens holocaustum ecce Samu-

end of offering the burnt offering, behold, Samuel came; and Saul went out to meet him, that he might salute him.

hel ueniebat et egressus est Saul obuiam ei ut salutaret eum

13:11 And Samuel said, What hast thou done? And Saul 13:11 locutusque est ad eum Samuhel quid fecisti respondit

said, Because I saw that the people were scattered from me, and that thou camest not within the days appointed, and that the Philistines gathered themselves together at Michmash;

Saul quia uidi quod dilaberetur populus a me et tu non ueneras iuxta placitos dies porro Philisthim congregati fuerant in

13:12 Therefore said I, The Philistines will come down Machmas 13:12 dixi nunc descendent Philisthim ad me in now upon me to Gilgal, and I have not made supplication unto the LORD: I forced myself therefore, and offered a

burnt offering. Galgala et faciem Domini non placaui necessitate conpulsus 13:13 And Samuel said to Saul, Thou hast done fool- obtuli holocaustum 13:13 dixitque Samuhel ad Saul stulte

ishly: thou hast not kept the commandment of the LORD thy God, which he commanded thee: for now would the

LORD have established thy kingdom upon Israel for ever. egisti nec custodisti mandata Domini Dei tui quae praece-

pit tibi quod si non fecisses iam nunc praeparasset Domi-

13:14 But now thy kingdom shall not continue: the LORD nus regnum tuum super Israhel in sempiternum 13:14 sed

hath sought him a man after his own heart, and the LORD hath commanded him to be captain over his people, because thou hast not kept that which the LORD commanded thee.

nequaquam regnum tuum ultra consurget quaesiuit sibi Dominus uirum iuxta cor suum et praecepit ei Dominus ut esset dux super populum suum eo quod non seruaueris quae

13:15 And Samuel arose, and gat him up from Gilgal unto praecepit Dominus 13:15 surrexit autem Samuhel et ascen-

Gibeah of Benjamin. And Saul numbered the people that were present with him, about six hundred men.

dit de Galgalis in Gabaa Beniamin et recensuit Saul popu-

13:16 And Saul, and Jonathan his son, and the people that lum qui inuenti fuerant cum eo quasi sescentos uiros 13:16

were present with them, abode in Gibeah of Benjamin: but the Philistines encamped in Michmash.

et Saul et Ionathan filius eius populusque qui inuentus fuerat cum eis erat in Gabaa Beniamin porro Philisthim con-

13:17 And the spoilers came out of the camp of the sederant in Machmas 13:17 et egressi sunt ad praedandum

Philistines in three companies: one company turned unto the way that leadeth to Ophrah, unto the land of Shual:

de castris Philisthim tres cunei unus cuneus pergebat contra

13:18 And another company turned the way to Bethhoron: uiam Ephra ad terram Saul 13:18 porro alius ingredieba-

and another company turned to the way of the border that looketh to the valley of Zeboim toward the wilderness.

tur per uiam Bethoron tertius autem uerterat se ad iter ter-

13:19 Now there was no smith found throughout all the mini inminentis ualli Seboim contra desertum 13:19 porro

land of Israel: for the Philistines said, Lest the Hebrews make them swords or spears:

faber ferrarius non inueniebatur in omni terra Israhel cauerant enim Philisthim ne forte facerent Hebraei gladium aut

13:20 But all the Israelites went down to the Philistines, to lanceam 13:20 descendebat ergo omnis Israhel ad Philistsharpen every man his share, and his coulter, and his axe, and his mattock.

him ut exacueret unusquisque uomerem suum et ligonem et

13:21 Yet they had a file for the mattocks, and for the coul- securim et sarculum 13:21 retunsae itaque erant acies uo-

ters, and for the forks, and for the axes, and to sharpen the goads.

merum et ligonum et tridentum et securium usque ad stimu-

13:22 So it came to pass in the day of battle, that there lum corrigendum 13:22 cumque uenisset dies proelii non

was neither sword nor spear found in the hand of any of the people that were with Saul and Jonathan: but with Saul

and with Jonathan his son was there found. est inuentus ensis et lancea in manu totius populi qui erat

Liber I Samuelis

cum Saul et cum Ionathan excepto Saul et Ionathan filio eius 13:23 egressa est autem statio Philisthim ut transcenderet in 13:23 And the garrison of the Philistines went out to the

passage of Michmash.

Machmas

14:1 et accidit quadam die ut diceret Ionathan filius Saul 14:1 Now it came to pass upon a day, that Jonathan the

son of Saul said unto the young man that bare his armour, Come, and let us go over to the Philistines’ garrison, that

ad adulescentem armigerum suum ueni et transeamus ad sta- is on the other side. But he told not his father. tionem Philisthim quae est trans locum illum patri autem

suo hoc ipsum non indicauit 14:2 porro Saul morabatur in 14:2 And Saul tarried in the uttermost part of Gibeah under a pomegranate tree which is in Migron: and the people that were with him were about six hundred men;

extrema parte Gabaa sub malogranato quae erat in Magron et erat populus cum eo quasi sescentorum uirorum 14:3 14:3 And Ahiah, the son of Ahitub, Ichabod’s brother,

the son of Phinehas, the son of Eli, the LORD’s priest in Shiloh, wearing an ephod. And the people knew not that

et Ahias filius Achitob fratris Ichabod filii Finees qui ortus Jonathan was gone. fuerat ex Heli sacerdote Domini in Silo portabat ephod sed et populus ignorabat quod isset Ionathan 14:4 erant autem 14:4 And between the passages, by which Jonathan sought

to go over unto the Philistines’ garrison, there was a sharp rock on the one side, and a sharp rock on the other side: and the name of the one was Bozez, and the name of the other Seneh.

inter ascensus per quos nitebatur Ionathan transire ad stationem Philisthinorum eminentes petrae ex utraque parte et quasi in modum dentium scopuli hinc inde praerupti nomen uni Boses et nomen alteri Sene 14:5 unus scopulus promi- 14:5 The forefront of the one was situate northward over

against Michmash, and the other southward over against Gibeah.

nens ad aquilonem ex aduerso Machmas et alter a meridie contra Gabaa 14:6 dixit autem Ionathan ad adulescentem 14:6 And Jonathan said to the young man that bare his

armour, Come, and let us go over unto the garrison of these uncircumcised: it may be that the LORD will work for us: for there is no restraint to the LORD to save by many or by few.

armigerum suum ueni transeamus ad stationem incircumcisorum horum si forte faciat Dominus pro nobis quia non est Domino difficile saluare uel in multitudine uel in paucis 14:7 dixitque ei armiger suus fac omnia quae placent animo 14:7 And his armourbearer said unto him, Do all that is in

thine heart: turn thee; behold, I am with thee according to thy heart.

tuo perge quo cupis ero tecum ubicumque uolueris 14:8 et

14:8 Then said Jonathan, Behold, we will pass over unto these men, and we will discover ourselves unto them.

ait Ionathan ecce nos transimus ad uiros istos cumque apparuerimus eis 14:9 si taliter locuti fuerint ad nos manete 14:9 If they say thus unto us, Tarry until we come to you;

then we will stand still in our place, and will not go up unto them.

donec ueniamus ad uos stemus in loco nostro nec ascendamus ad eos 14:10 si autem dixerint ascendite ad nos ascen- 14:10 But if they say thus, Come up unto us; then we will

go up: for the LORD hath delivered them into our hand: and this shall be a sign unto us.

damus quia tradidit eos Dominus in manibus nostris hoc erit nobis signum 14:11 apparuit igitur uterque stationi Philis- 14:11 And both of them discovered themselves unto the

garrison of the Philistines: and the Philistines said, Behold, the Hebrews come forth out of the holes where they

thinorum dixeruntque Philisthim en Hebraei egrediuntur de had hid themselves. cauernis in quibus absconditi fuerant 14:12 et locuti sunt 14:12 And the men of the garrison answered Jonathan

and his armourbearer, and said, Come up to us, and we will shew you a thing. And Jonathan said unto his armourbearer, Come up after me: for the LORD hath delivered them into the hand of Israel.

uiri de statione ad Ionathan et ad armigerum eius dixeruntque ascendite ad nos et ostendimus uobis rem et ait Ionathan

Biblia Sacra Vulgata

ad armigerum suum ascendamus sequere me tradidit enim

14:13 And Jonathan climbed up upon his hands and upon eos Dominus in manu Israhel 14:13 ascendit autem Iona-

his feet, and his armourbearer after him: and they fell before Jonathan; and his armourbearer slew after him.

than reptans manibus et pedibus et armiger eius post eum itaque alii cadebant ante Ionathan alios armiger eius inter-

14:14 And that first slaughter, which Jonathan and his ar- ficiebat sequens eum 14:14 et facta est plaga prima quam mourbearer made, was about twenty men, within as it were an half acre of land, which a yoke of oxen might plow.

percussit Ionathan et armiger eius quasi uiginti uirorum in media parte iugeri quam par boum in die arare consueuit

14:15 And there was trembling in the host, in the field, and 14:15 et factum est miraculum in castris per agros sed et om-

among all the people: the garrison, and the spoilers, they also trembled, and the earth quaked: so it was a very great

trembling. nis populus stationis eorum qui ierant ad praedandum ob-

stipuit et conturbata est terra et accidit quasi miraculum a

14:16 And the watchmen of Saul in Gibeah of Benjamin Deo 14:16 et respexerunt speculatores Saul qui erant in Ga-

looked; and, behold, the multitude melted away, and they went on beating down one another.

baa Beniamin et ecce multitudo prostrata et huc illucque dif-

14:17 Then said Saul unto the people that were with him, fugiens 14:17 et ait Saul populo qui erat cum eo requirite et

Number now, and see who is gone from us. And when they had numbered, behold, Jonathan and his armourbearer

were not there. uidete quis abierit ex nobis cumque requisissent reppertum 14:18 And Saul said unto Ahiah, Bring hither the ark of est non adesse Ionathan et armigerum eius 14:18 et ait Saul

God. For the ark of God was at that time with the children of Israel.

ad Ahiam adplica arcam Dei erat enim ibi arca Dei in die

14:19 And it came to pass, while Saul talked unto the illa cum filiis Israhel 14:19 cumque loqueretur Saul ad sa-

priest, that the noise that was in the host of the Philistines went on and increased: and Saul said unto the priest, With-

draw thine hand. cerdotem tumultus magnus exortus est in castris Philisthino-

rum crescebatque paulatim et clarius reboabat et ait Saul ad

14:20 And Saul and all the people that were with him as- sacerdotem contrahe manum tuam 14:20 conclamauit ergo

sembled themselves, and they came to the battle: and, behold, every man’s sword was against his fellow, and there

was a very great discomfiture. Saul et omnis populus qui erat cum eo et uenerunt usque ad

locum certaminis et ecce uersus fuerat gladius uniuscuius-

14:21 Moreover the Hebrews that were with the Philistines que ad proximum suum et caedes magna nimis 14:21 sed before that time, which went up with them into the camp from the country round about, even they also turned to be

with the Israelites that were with Saul and Jonathan. et Hebraei qui fuerant cum Philisthim heri et nudius tertius

ascenderantque cum eis in castris reuersi sunt ut essent cum

14:22 Likewise all the men of Israel which had hid Israhele qui erant cum Saul et Ionathan 14:22 omnes quo-

themselves in mount Ephraim, when they heard that the Philistines fled, even they also followed hard after them in

the battle. que Israhelitae qui se absconderant in monte Ephraim au-

dientes quod fugissent Philisthim sociauerunt se cum suis in

14:23 So the LORD saved Israel that day: and the battle proelio 14:23 et saluauit Dominus in die illa Israhel pugna

passed over unto Bethaven.

14:24 And the men of Israel were distressed that day: for autem peruenit usque Bethauen 14:24 et uir Israhel sociatus

Saul had adjured the people, saying, Cursed be the man that eateth any food until evening, that I may be avenged

on mine enemies. So none of the people tasted any food. sibi est in die illa adiurauit autem Saul populum dicens maledictus uir qui comederit panem usque ad uesperam donec ulciscar de inimicis meis et non manducauit uniuersus po-

Liber I Samuelis

pulus panem 14:25 omneque terrae uulgus uenit in saltum 14:25 And all they of the land came to a wood; and there

was honey upon the ground.

in quo erat mel super faciem agri 14:26 ingressus est ita- 14:26 And when the people were come into the wood, be-

hold, the honey dropped; but no man put his hand to his mouth: for the people feared the oath.

que populus saltum et apparuit fluens mel nullusque adplicuit manum ad os suum timebat enim populus iuramentum 14:27 porro Ionathan non audierat cum adiuraret pater eius 14:27 But Jonathan heard not when his father charged the

people with the oath: wherefore he put forth the end of the rod that was in his hand, and dipped it in an honeycomb, and put his hand to his mouth; and his eyes were enlightened.

populum extenditque summitatem uirgae quam habebat in manu et intinxit in fauo mellis et conuertit manum suam ad os suum et inluminati sunt oculi eius 14:28 respondensque 14:28 Then answered one of the people, and said, Thy

father straitly charged the people with an oath, saying, Cursed be the man that eateth any food this day. And the

unus de populo ait iureiurando constrinxit pater tuus popu- people were faint. lum dicens maledictus qui comederit panem hodie defecerat autem populus 14:29 dixitque Ionathan turbauit pater meus 14:29 Then said Jonathan, My father hath troubled the

land: see, I pray you, how mine eyes have been enlightened, because I tasted a little of this honey.

terram uidistis ipsi quia inluminati sunt oculi mei eo quod gustauerim paululum de melle isto 14:30 quanto magis si 14:30 How much more, if haply the people had eaten

freely to day of the spoil of their enemies which they found? for had there not been now a much greater slaugh-

comedisset populus de praeda inimicorum suorum quam re- ter among the Philistines? pperit nonne maior facta fuisset plaga in Philisthim 14:31 14:31 And they smote the Philistines that day from Mich-

mash to Aijalon: and the people were very faint.

percusserunt ergo in die illa Philistheos a Machmis usque in Ahialon defatigatus est autem populus nimis 14:32 et uer- 14:32 And the people flew upon the spoil, and took sheep,

and oxen, and calves, and slew them on the ground: and the people did eat them with the blood.

sus ad praedam tulit oues et boues et uitulos et mactauerunt in terra comeditque populus cum sanguine 14:33 nuntiaue- 14:33 Then they told Saul, saying, Behold, the people sin

against the LORD, in that they eat with the blood. And he said, Ye have transgressed: roll a great stone unto me this

runt autem Saul dicentes quod populus peccasset Domino day. comedens cum sanguine qui ait praeuaricati estis uoluite ad me iam nunc saxum grande 14:34 et dixit Saul dispergimini 14:34 And Saul said, Disperse yourselves among the peo-

ple, and say unto them, Bring me hither every man his ox, and every man his sheep, and slay them here, and eat; and sin not against the LORD in eating with the blood. And all the people brought every man his ox with him that night, and slew them there.

in uulgus et dicite eis ut adducat ad me unusquisque bouem suum et arietem et occidite super istud et uescimini et non peccabitis Domino comedentes cum sanguine adduxit itaque omnis populus unusquisque bouem in manu sua usque ad noctem et occiderunt ibi 14:35 aedificauit autem Saul al- 14:35 And Saul built an altar unto the LORD: the same

was the first altar that he built unto the LORD.

tare Domini tuncque primum coepit aedificare altare Domini 14:36 et dixit Saul inruamus super Philisthim nocte et uaste- 14:36 And Saul said, Let us go down after the Philistines

by night, and spoil them until the morning light, and let us not leave a man of them. And they said, Do whatsoever seemeth good unto thee. Then said the priest, Let us draw near hither unto God.

mus eos usque dum inlucescat mane nec relinquamus de eis uirum dixitque populus omne quod bonum uidetur in oculis tuis fac et ait sacerdos accedamus huc ad Deum 14:37 et 14:37 And Saul asked counsel of God, Shall I go down

after the Philistines? wilt thou deliver them into the hand of Israel? But he answered him not that day.

consuluit Saul Deum num persequar Philisthim si trades eos

Biblia Sacra Vulgata

14:38 And Saul said, Draw ye near hither, all the chief of in manu Israhel et non respondit ei in die illa 14:38 dixit-

the people: and know and see wherein this sin hath been this day.

que Saul adplicate huc uniuersos angulos populi et scitote et

14:39 For, as the LORD liveth, which saveth Israel, though uidete per quem acciderit peccatum hoc hodie 14:39 uiuit it be in Jonathan my son, he shall surely die. But there was not a man among all the people that answered him.

Dominus saluator Israhel quia si per Ionathan filium meum factum est absque retractatione morietur ad quod nullus con-

14:40 Then said he unto all Israel, Be ye on one side, and tradixit ei de omni populo 14:40 et ait ad uniuersum Israhel

I and Jonathan my son will be on the other side. And the people said unto Saul, Do what seemeth good unto thee.

separamini uos in partem unam et ego cum Ionathan filio meo ero in parte una respondit populus ad Saul quod bo-

14:41 Therefore Saul said unto the LORD God of Israel, num uidetur in oculis tuis fac 14:41 et dixit Saul ad Domi-

Give a perfect lot. And Saul and Jonathan were taken: but the people escaped.

num Deum Israhel da indicium et deprehensus est Ionathan

14:42 And Saul said, Cast lots between me and Jonathan et Saul populus autem exiuit 14:42 et ait Saul mittite sor-

my son. And Jonathan was taken.

tem inter me et inter Ionathan filium meum et captus est

14:43 Then Saul said to Jonathan, Tell me what thou hast Ionathan 14:43 dixit autem Saul ad Ionathan indica mihi

done. And Jonathan told him, and said, I did but taste a little honey with the end of the rod that was in mine hand,

and, lo, I must die. quid feceris et indicauit ei Ionathan et ait gustans gustaui in

summitate uirgae quae erat in manu mea paululum mellis

14:44 And Saul answered, God do so and more also: for et ecce ego morior 14:44 et ait Saul haec faciat mihi Deus

thou shalt surely die, Jonathan.

14:45 And the people said unto Saul, Shall Jonathan die, et haec addat quia morte morieris Ionathan 14:45 dixitque

who hath wrought this great salvation in Israel? God forbid: as the LORD liveth, there shall not one hair of his head fall to the ground; for he hath wrought with God this day. So the people rescued Jonathan, that he died not.

populus ad Saul ergone Ionathan morietur qui fecit salutem hanc magnam in Israhel hoc nefas est uiuit Dominus si ceciderit capillus de capite eius in terram quia cum Deo operatus est hodie liberauit ergo populus Ionathan ut non mo-

14:46 Then Saul went up from following the Philistines: reretur 14:46 recessitque Saul nec persecutus est Philisthim

and the Philistines went to their own place.

14:47 So Saul took the kingdom over Israel, and fought porro Philisthim abierunt in loca sua 14:47 at Saul confir-

against all his enemies on every side, against Moab, and against the children of Ammon, and against Edom, and against the kings of Zobah, and against the Philistines: and whithersoever he turned himself, he vexed them.

mato regno super Israhel pugnabat per circuitum aduersum omnes inimicos eius contra Moab et filios Ammon et Edom et reges Suba et Philistheos et quocumque se uerterat su-

14:48 And he gathered an host, and smote the Amalekites, perabat 14:48 congregatoque exercitu percussit Amalech et

and delivered Israel out of the hands of them that spoiled them.

eruit Israhel de manu uastatorum eius 14:49 fuerunt autem

14:49 Now the sons of Saul were Jonathan, and Ishui, and Melchishua: and the names of his two daughters were these; the name of the firstborn Merab, and the name of the younger Michal:

filii Saul Ionathan et Iesui et Melchisua nomina duarum filiarum eius nomen primogenitae Merob et nomen minoris

14:50 And the name of Saul’s wife was Ahinoam, the Michol 14:50 et nomen uxoris Saul Ahinoem filia Ahimaas

daughter of Ahimaaz: and the name of the captain of his host was Abner, the son of Ner, Saul’s uncle.

et nomina principum militiae eius Abner filius Ner patruelis

14:51 And Kish was the father of Saul; and Ner the father Saul 14:51 Cis fuerat pater Saul et Ner pater Abner filius

of Abner was the son of Abiel.

Liber I Samuelis

Abihel 14:52 erat autem bellum potens aduersum Philist- 14:52 And there was sore war against the Philistines all

the days of Saul: and when Saul saw any strong man, or any valiant man, he took him unto him.

heos omnibus diebus Saul nam quemcumque uiderat Saul uirum fortem et aptum ad proelium sociabat eum sibi

15:1 et dixit Samuhel ad Saul me misit Dominus ut un- 15:1 Samuel also said unto Saul, The LORD sent me to

anoint thee to be king over his people, over Israel: now therefore hearken thou unto the voice of the words of the

guerem te in regem super populum eius Israhel nunc ergo LORD.

audi uocem Domini 15:2 haec dicit Dominus exercituum 15:2 Thus saith the LORD of hosts, I remember that which Amalek did to Israel, how he laid wait for him in the way, when he came up from Egypt.

recensui quaecumque fecit Amalech Israheli quomodo restitit ei in uia cum ascenderet de Aegypto 15:3 nunc igitur 15:3 Now go and smite Amalek, and utterly destroy all

that they have, and spare them not; but slay both man and woman, infant and suckling, ox and sheep, camel and ass.

uade et percute Amalech et demolire uniuersa eius non parcas ei sed interfice a uiro usque ad mulierem et paruulum

atque lactantem bouem et ouem camelum et asinum 15:4 15:4 And Saul gathered the people together, and numbered them in Telaim, two hundred thousand footmen, and ten thousand men of Judah.

praecepit itaque Saul populo et recensuit eos quasi agnos ducenta milia peditum et decem milia uirorum Iuda 15:5 15:5 And Saul came to a city of Amalek, and laid wait in

the valley.

cumque uenisset Saul usque ad ciuitatem Amalech tetendit insidias in torrente 15:6 dixitque Saul Cineo abite recedite 15:6 And Saul said unto the Kenites, Go, depart, get you

down from among the Amalekites, lest I destroy you with them: for ye shewed kindness to all the children of Israel, when they came up out of Egypt. So the Kenites departed from among the Amalekites.

atque descendite ab Amalech ne forte inuoluam te cum eo tu enim fecisti misericordiam cum omnibus filiis Israhel cum ascenderent de Aegypto et recessit Cineus de medio Amalech 15:7 percussitque Saul Amalech ab Euila donec uenias 15:7 And Saul smote the Amalekites from Havilah until

thou comest to Shur, that is over against Egypt.

Sur quae est e regione Aegypti 15:8 et adprehendit Agag 15:8 And he took Agag the king of the Amalekites alive,

and utterly destroyed all the people with the edge of the sword.

regem Amalech uiuum omne autem uulgus interfecit in ore gladii 15:9 et pepercit Saul et populus Agag et optimis gre- 15:9 But Saul and the people spared Agag, and the best

of the sheep, and of the oxen, and of the fatlings, and the lambs, and all that was good, and would not utterly destroy them: but every thing that was vile and refuse, that they destroyed utterly.

gibus ouium et armentorum et uestibus et arietibus et uniuersis quae pulchra erant nec uoluerunt disperdere ea quicquid uero uile fuit et reprobum hoc demoliti sunt 15:10 factum 15:10 Then came the word of the LORD unto Samuel, say-

ing,

est autem uerbum Domini ad Samuhel dicens 15:11 pae- 15:11 It repenteth me that I have set up Saul to be king: for he is turned back from following me, and hath not performed my commandments. And it grieved Samuel; and

nitet me quod constituerim Saul regem quia dereliquit me he cried unto the LORD all night. et uerba mea opere non impleuit contristatusque est Samuhel et clamauit ad Dominum tota nocte 15:12 cumque de 15:12 And when Samuel rose early to meet Saul in the

morning, it was told Samuel, saying, Saul came to Carmel, and, behold, he set him up a place, and is gone about, and

nocte surrexisset Samuhel ut iret ad Saul mane nuntiatum passed on, and gone down to Gilgal. est Samuheli eo quod uenisset Saul in Carmelum et erexisset sibi fornicem triumphalem et reuersus transisset descendissetque in Galgala uenit ergo Samuhel ad Saul et 15:13 15:13 And Samuel came to Saul: and Saul said unto him,

Blessed be thou of the LORD: I have performed the commandment of the LORD.

Biblia Sacra Vulgata

dixit ei Saul benedictus tu Domino impleui uerbum Domini

15:14 And Samuel said, What meaneth then this bleating 15:14 dixitque Samuhel et quae est haec uox gregum quae re-

of the sheep in mine ears, and the lowing of the oxen which I hear?

sonat in auribus meis et armentorum quam ego audio 15:15

15:15 And Saul said, They have brought them from the Amalekites: for the people spared the best of the sheep and of the oxen, to sacrifice unto the LORD thy God; and the rest we have utterly destroyed.

et ait Saul de Amalech adduxerunt ea pepercit enim populus melioribus ouibus et armentis ut immolarentur Domino

15:16 Then Samuel said unto Saul, Stay, and I will tell Deo tuo reliqua uero occidimus 15:16 dixit autem Samuhel

thee what the LORD hath said to me this night. And he said unto him, Say on.

ad Saul sine me et indicabo tibi quae locutus sit Dominus

15:17 And Samuel said, When thou wast little in thine own ad me nocte dixitque ei loquere 15:17 et ait Samuhel nonne

sight, wast thou not made the head of the tribes of Israel, and the LORD anointed thee king over Israel?

cum paruulus esses in oculis tuis caput in tribubus Israhel

15:18 And the LORD sent thee on a journey, and said, Go factus es unxitque te Dominus regem super Israhel 15:18

and utterly destroy the sinners the Amalekites, and fight against them until they be consumed.

et misit te Dominus in uia et ait uade et interfice peccatores Amalech et pugnabis contra eos usque ad internicionem

15:19 Wherefore then didst thou not obey the voice of the eorum 15:19 quare ergo non audisti uocem Domini sed uer-LORD, but didst fly upon the spoil, and didst evil in the sight of the LORD?

sus ad praedam es et fecisti malum in oculis Domini 15:20

15:20 And Saul said unto Samuel, Yea, I have obeyed the voice of the LORD, and have gone the way which the LORD sent me, and have brought Agag the king of Amalek, and have utterly destroyed the Amalekites.

et ait Saul ad Samuhelem immo audiui uocem Domini et ambulaui in uia per quam misit me Dominus et adduxi Agag

15:21 But the people took of the spoil, sheep and oxen, regem Amalech et Amalech interfeci 15:21 tulit autem po-

the chief of the things which should have been utterly destroyed, to sacrifice unto the LORD thy God in Gilgal.

pulus de praeda oues et boues primitias eorum quae caesa

15:22 And Samuel said, Hath the LORD as great delight sunt ut immolet Domino Deo suo in Galgalis 15:22 et ait

in burnt offerings and sacrifices, as in obeying the voice of the LORD? Behold, to obey is better than sacrifice, and to

hearken than the fat of rams. Samuhel numquid uult Dominus holocausta aut uictimas et non potius ut oboediatur uoci Domini melior est enim oboedientia quam uictimae et auscultare magis quam offerre ad-

15:23 For rebellion is as the sin of witchcraft, and stub- ipem arietum 15:23 quoniam quasi peccatum ariolandi est bornness is as iniquity and idolatry. Because thou hast rejected the word of the LORD, he hath also rejected thee

from being king. repugnare et quasi scelus idolatriae nolle adquiescere pro eo

ergo quod abiecisti sermonem Domini abiecit te ne sis rex

15:24 And Saul said unto Samuel, I have sinned: for I 15:24 dixitque Saul ad Samuhel peccaui quia praeuaricatus

have transgressed the commandment of the LORD, and thy words: because I feared the people, and obeyed their

voice. sum sermonem Domini et uerba tua timens populum et ob-

15:25 Now therefore, I pray thee, pardon my sin, and turn oediens uoci eorum 15:25 sed nunc porta quaeso peccatum again with me, that I may worship the LORD. 15:26 And Samuel said unto Saul, I will not return with meum et reuertere mecum ut adorem Dominum 15:26 et thee: for thou hast rejected the word of the LORD, and the LORD hath rejected thee from being king over Israel.

ait Samuhel ad Saul non reuertar tecum quia proiecisti sermonem Domini et proiecit te Dominus ne sis rex super Is-

15:27 And as Samuel turned about to go away, he laid hold rahel 15:27 et conuersus est Samuhel ut abiret ille autem

upon the skirt of his mantle, and it rent.

15:28 And Samuel said unto him, The LORD hath rent the adprehendit summitatem pallii eius quae et scissa est 15:28

kingdom of Israel from thee this day, and hath given it to a neighbour of thine, that is better than thou.

Liber I Samuelis

et ait ad eum Samuhel scidit Dominus regnum Israhel a te hodie et tradidit illud proximo tuo meliori te 15:29 porro 15:29 And also the Strength of Israel will not lie nor re-

pent: for he is not a man, that he should repent.

Triumphator in Israhel non parcet et paenitudine non flectetur neque enim homo est ut agat paenitentiam 15:30 at 15:30 Then he said, I have sinned: yet honour me now, I

pray thee, before the elders of my people, and before Israel, and turn again with me, that I may worship the LORD

ille ait peccaui sed nunc honora me coram senibus populi thy God. mei et coram Israhel et reuertere mecum ut adorem Dominum Deum tuum 15:31 reuersus ergo Samuhel secutus est 15:31 So Samuel turned again after Saul; and Saul wor-

shipped the LORD.

Saulem et adorauit Saul Dominum 15:32 dixitque Samu- 15:32 Then said Samuel, Bring ye hither to me Agag the

king of the Amalekites. And Agag came unto him delicately. And Agag said, Surely the bitterness of death is

hel adducite ad me Agag regem Amalech et oblatus est ei past. Agag pinguissimus et dixit Agag sicine separat amara mors 15:33 et ait Samuhel sicut fecit absque liberis mulieres gla- 15:33 And Samuel said, As the sword hath made women

childless, so shall thy mother be childless among women. And Samuel hewed Agag in pieces before the LORD in

dius tuus sic absque liberis erit inter mulieres mater tua et Gilgal. in frusta concidit Samuhel Agag coram Domino in Galgalis 15:34 abiit autem Samuhel in Ramatha Saul uero ascendit in 15:34 Then Samuel went to Ramah; and Saul went up to

his house to Gibeah of Saul.

domum suam in Gabaath 15:35 et non uidit Samuhel ultra 15:35 And Samuel came no more to see Saul until the day

of his death: nevertheless Samuel mourned for Saul: and the LORD repented that he had made Saul king over Israel.

Saul usque ad diem mortis suae uerumtamen lugebat Samuhel Saul quoniam Dominum paenitebat quod constituisset regem Saul super Israhel

16:1 dixitque Dominus ad Samuhel usquequo tu luges Saul 16:1 And the LORD said unto Samuel, How long wilt thou

mourn for Saul, seeing I have rejected him from reigning over Israel? fill thine horn with oil, and go, I will send thee to Jesse the Bethlehemite: for I have provided me a king among his sons.

cum ego proiecerim eum ne regnet super Israhel imple cornu tuum oleo et ueni ut mittam te ad Isai Bethleemitem prouidi enim in filiis eius mihi regem 16:2 et ait Samuhel quomodo 16:2 And Samuel said, How can I go? if Saul hear it, he

will kill me. And the LORD said, Take an heifer with thee, and say, I am come to sacrifice to the LORD.

uadam audiet enim Saul et interficiet me et ait Dominus uitulum de armento tolles in manu tua et dices ad immolandum Domino ueni 16:3 et uocabis Isai ad uictimam et ego osten- 16:3 And call Jesse to the sacrifice, and I will shew thee

what thou shalt do: and thou shalt anoint unto me him whom I name unto thee.

dam tibi quid facias et ungues quemcumque monstrauero tibi 16:4 fecit ergo Samuhel sicut locutus est ei Dominus uenitque 16:4 And Samuel did that which the LORD spake, and

came to Bethlehem. And the elders of the town trembled at his coming, and said, Comest thou peaceably?

in Bethleem et admirati sunt seniores ciuitatis occurrentes ei dixeruntque pacificus ingressus tuus 16:5 et ait pacificus ad 16:5 And he said, Peaceably: I am come to sacrifice unto

the LORD: sanctify yourselves, and come with me to the sacrifice. And he sanctified Jesse and his sons, and called

immolandum Domino ueni sanctificamini et uenite mecum them to the sacrifice. ut immolem sanctificauit ergo Isai et filios eius et uocauit eos ad sacrificium 16:6 cumque ingressi essent uidit Heliab et 16:6 And it came to pass, when they were come, that he

looked on Eliab, and said, Surely the LORD’s anointed is before him.

ait num coram Domino est christus eius 16:7 et dixit Domi-

16:7 But the LORD said unto Samuel, Look not on his countenance, or on the height of his stature; because I have refused him: for the LORD seeth not as man seeth; for man looketh on the outward appearance, but the LORD looketh on the heart.

CDLXXIX

Biblia Sacra Vulgata

nus ad Samuhel ne respicias uultum eius neque altitudinem staturae eius quoniam abieci eum nec iuxta intuitum hominis iudico homo enim uidet ea quae parent Dominus autem

16:8 Then Jesse called Abinadab, and made him pass be- intuetur cor 16:8 et uocauit Isai Abinadab et adduxit eum

fore Samuel. And he said, Neither hath the LORD chosen this.

coram Samuhel qui dixit nec hunc elegit Dominus 16:9 ad-

16:9 Then Jesse made Shammah to pass by. And he said, Neither hath the LORD chosen this.

duxit autem Isai Samma de quo ait etiam hunc non elegit

16:10 Again, Jesse made seven of his sons to pass before Dominus 16:10 adduxit itaque Isai septem filios suos coram

Samuel. And Samuel said unto Jesse, The LORD hath not chosen these.

Samuhel et ait Samuhel ad Isai non elegit Dominus ex istis

16:11 And Samuel said unto Jesse, Are here all thy chil- 16:11 dixitque Samuhel ad Isai numquid iam conpleti sunt fi-

dren? And he said, There remaineth yet the youngest, and, behold, he keepeth the sheep. And Samuel said unto Jesse, Send and fetch him: for we will not sit down till he come hither.

lii qui respondit adhuc reliquus est paruulus et pascit oues et ait Samuhel ad Isai mitte et adduc eum nec enim discumbe-

16:12 And he sent, and brought him in. Now he was ruddy, mus priusquam ille huc uenerit 16:12 misit ergo et adduxit

and withal of a beautiful countenance, and goodly to look to. And the LORD said, Arise, anoint him: for this is he.

eum erat autem rufus et pulcher aspectu decoraque facie et

16:13 Then Samuel took the horn of oil, and anointed him ait Dominus surge ungue eum ipse est enim 16:13 tulit igi-

in the midst of his brethren: and the Spirit of the LORD came upon David from that day forward. So Samuel rose

up, and went to Ramah. tur Samuhel cornu olei et unxit eum in medio fratrum eius et

directus est spiritus Domini in Dauid a die illa et in reliquum

16:14 But the Spirit of the LORD departed from Saul, and surgensque Samuhel abiit in Ramatha 16:14 spiritus autem

an evil spirit from the LORD troubled him.

Domini recessit a Saul et exagitabat eum spiritus nequam a

16:15 And Saul’s servants said unto him, Behold now, an Domino 16:15 dixeruntque serui Saul ad eum ecce spiritus

evil spirit from God troubleth thee.

16:16 Let our lord now command thy servants, which are Dei malus exagitat te 16:16 iubeat dominus noster et serui

before thee, to seek out a man, who is a cunning player on an harp: and it shall come to pass, when the evil spirit from God is upon thee, that he shall play with his hand, and thou shalt be well.

tui qui coram te sunt quaerant hominem scientem psallere cithara ut quando arripuerit te spiritus Dei malus psallat manu

16:17 And Saul said unto his servants, Provide me now a sua et leuius feras 16:17 et ait Saul ad seruos suos prouidete

man that can play well, and bring him to me.

16:18 Then answered one of the servants, and said, Be- mihi aliquem bene psallentem et adducite eum ad me 16:18

hold, I have seen a son of Jesse the Bethlehemite, that is cunning in playing, and a mighty valiant man, and a man of war, and prudent in matters, and a comely person, and the LORD is with him.

et respondens unus de pueris ait ecce uidi filium Isai Bethleemitem scientem psallere et fortissimum robore et uirum bellicosum et prudentem in uerbis et uirum pulchrum et Do-

16:19 Wherefore Saul sent messengers unto Jesse, and minus est cum eo 16:19 misit ergo Saul nuntios ad Isai di-

said, Send me David thy son, which is with the sheep. 16:20 And Jesse took an ass laden with bread, and a bottle cens mitte ad me Dauid filium tuum qui est in pascuis 16:20 of wine, and a kid, and sent them by David his son unto Saul.

tulitque Isai asinum plenum panibus et lagoenam uini et hedum de capris unum et misit per manum Dauid filii sui Saul

16:21 And David came to Saul, and stood before him: and 16:21 et uenit Dauid ad Saul et stetit coram eo at ille dilexit

he loved him greatly; and he became his armourbearer.

16:22 And Saul sent to Jesse, saying, Let David, I pray eum nimis et factus est eius armiger 16:22 misitque Saul ad

thee, stand before me; for he hath found favour in my sight.

Liber I Samuelis

Isai dicens stet Dauid in conspectu meo inuenit enim gratiam in oculis meis 16:23 igitur quandocumque spiritus Dei 16:23 And it came to pass, when the evil spirit from God

was upon Saul, that David took an harp, and played with his hand: so Saul was refreshed, and was well, and the evil

arripiebat Saul tollebat Dauid citharam et percutiebat manu spirit departed from him. sua et refocilabatur Saul et leuius habebat recedebat enim ab eo spiritus malus

17:1 congregantes uero Philisthim agmina sua in proelium 17:1 Now the Philistines gathered together their armies

to battle, and were gathered together at Shochoh, which belongeth to Judah, and pitched between Shochoh and

conuenerunt in Soccho Iudae et castrametati sunt inter Soc- Azekah, in Ephesdammim. cho et Azeca in finibus Dommim 17:2 porro Saul et uiri 17:2 And Saul and the men of Israel were gathered to-

gether, and pitched by the valley of Elah, and set the battle in array against the Philistines.

Israhel congregati uenerunt in ualle Terebinthi et direxerunt aciem ad pugnandum contra Philisthim 17:3 et Philisthim 17:3 And the Philistines stood on a mountain on the one

side, and Israel stood on a mountain on the other side: and there was a valley between them.

stabant super montem ex hac parte et Israhel stabat super montem ex altera parte uallisque erat inter eos 17:4 et eg- 17:4 And there went out a champion out of the camp of

the Philistines, named Goliath, of Gath, whose height was six cubits and a span.

ressus est uir spurius de castris Philisthinorum nomine Goliath de Geth altitudinis sex cubitorum et palmo 17:5 et cas- 17:5 And he had an helmet of brass upon his head, and he

was armed with a coat of mail; and the weight of the coat was five thousand shekels of brass.

sis aerea super caput eius et lorica hamata induebatur porro pondus loricae eius quinque milia siclorum aeris 17:6 et 17:6 And he had greaves of brass upon his legs, and a

target of brass between his shoulders.

ocreas aereas habebat in cruribus et clypeus aereus tegebat umeros eius 17:7 hastile autem hastae eius erat quasi licia- 17:7 And the staff of his spear was like a weaver’s beam;

and his spear’s head weighed six hundred shekels of iron: and one bearing a shield went before him.

torium texentium ipsum autem ferrum hastae eius sescentos siclos habebat ferri et armiger eius antecedebat eum 17:8 17:8 And he stood and cried unto the armies of Israel, and

said unto them, Why are ye come out to set your battle in array? am not I a Philistine, and ye servants to Saul?

stansque clamabat aduersum falangas Israhel et dicebat eis choose you a man for you, and let him come down to me. quare uenitis parati ad proelium numquid ego non sum Philistheus et uos serui Saul eligite ex uobis uirum et descendat ad singulare certamen 17:9 si quiuerit pugnare mecum 17:9 If he be able to fight with me, and to kill me, then

will we be your servants: but if I prevail against him, and kill him, then shall ye be our servants, and serve us.

et percusserit me erimus uobis serui si autem ego praeualuero et percussero eum uos serui eritis et seruietis nobis 17:10 et aiebat Philistheus ego exprobraui agminibus Israhe- 17:10 And the Philistine said, I defy the armies of Israel

this day; give me a man, that we may fight together.

lis hodie date mihi uirum et ineat mecum singulare certamen 17:11 audiens autem Saul et omnes uiri israhelitae sermones 17:11 When Saul and all Israel heard those words of the

Philistine, they were dismayed, and greatly afraid.

Philisthei huiuscemodi stupebant et metuebant nimis 17:12 17:12 Now David was the son of that Ephrathite of Bethle-

hemjudah, whose name was Jesse; and he had eight sons: and the man went among men for an old man in the days

Dauid autem erat filius uiri ephrathei de quo supra dictum of Saul. est de Bethleem Iuda cui erat nomen Isai qui habebat octo filios et erat uir in diebus Saul senex et grandeuus inter ui-

Biblia Sacra Vulgata

17:13 And the three eldest sons of Jesse went and followed ros 17:13 abierunt autem tres filii eius maiores post Saul in

Saul to the battle: and the names of his three sons that went to the battle were Eliab the firstborn, and next unto

him Abinadab, and the third Shammah. proelium et nomina trium filiorum eius qui perrexerant ad

bellum Heliab primogenitus et secundus Abinadab tertius-

17:14 And David was the youngest: and the three eldest que Samma 17:14 Dauid autem erat minimus tribus ergo followed Saul.

17:15 But David went and returned from Saul to feed his maioribus secutis Saulem 17:15 abiit Dauid et reuersus est a

father’s sheep at Bethlehem.

17:16 And the Philistine drew near morning and evening, Saul ut pasceret gregem patris sui in Bethleem 17:16 proce-

and presented himself forty days.

debat uero Philistheus mane et uespere et stabat quadraginta

17:17 And Jesse said unto David his son, Take now for diebus 17:17 dixit autem Isai ad Dauid filium suum accipe

thy brethren an ephah of this parched corn, and these ten loaves, and run to the camp of thy brethren;

fratribus tuis oephi pulentae et decem panes istos et curre

17:18 And carry these ten cheeses unto the captain of their in castra ad fratres tuos 17:18 et decem formellas casei has

thousand, and look how thy brethren fare, and take their pledge.

deferes ad tribunum et fratres tuos uisitabis si recte agant et

17:19 Now Saul, and they, and all the men of Israel, were cum quibus ordinati sint disce 17:19 Saul autem et illi et

in the valley of Elah, fighting with the Philistines.

omnes filii Israhel in ualle Terebinthi pugnabant aduersum

17:20 And David rose up early in the morning, and left Philisthim 17:20 surrexit itaque Dauid mane et commenda-

the sheep with a keeper, and took, and went, as Jesse had commanded him; and he came to the trench, as the host

was going forth to the fight, and shouted for the battle. uit gregem custodi et onustus abiit sicut praeceperat ei Isai

et uenit ad locum Magala et ad exercitum qui egressus ad

17:21 For Israel and the Philistines had put the battle in pugnam uociferatus erat in certamine 17:21 direxerat enim

array, army against army.

aciem Israhel sed et Philisthim ex aduerso fuerant praepa-

17:22 And David left his carriage in the hand of the keeper rati 17:22 derelinquens ergo Dauid uasa quae adtulerat sub

of the carriage, and ran into the army, and came and saluted his brethren.

manu custodis ad sarcinas cucurrit ad locum certaminis et in-

17:23 And as he talked with them, behold, there came up terrogabat si omnia recte agerentur erga fratres suos 17:23

the champion, the Philistine of Gath, Goliath by name, out of the armies of the Philistines, and spake according to the

same words: and David heard them. cumque adhuc ille loqueretur eis apparuit uir ille spurius ascendens Goliath nomine Philistheus de Geth ex castris Philisthinorum et loquente eo haec eadem uerba audiuit Dauid

17:24 And all the men of Israel, when they saw the man, 17:24 omnes autem Israhelitae cum uidissent uirum fugerunt

fled from him, and were sore afraid.

17:25 And the men of Israel said, Have ye seen this man a facie eius timentes eum ualde 17:25 et dixit unus quis-

that is come up? surely to defy Israel is he come up: and it shall be, that the man who killeth him, the king will enrich him with great riches, and will give him his daughter, and make his father’s house free in Israel.

piam de Israhel num uidisti uirum hunc qui ascendit ad exprobrandum enim Israheli ascendit uirum ergo qui percusserit eum ditabit rex diuitiis magnis et filiam suam dabit ei et

17:26 And David spake to the men that stood by him, say- domum patris eius faciet absque tributo in Israhel 17:26 et

ing, What shall be done to the man that killeth this Philistine, and taketh away the reproach from Israel? for who is this uncircumcised Philistine, that he should defy the armies of the living God?

ait Dauid ad uiros qui stabant secum dicens quid dabitur uiro qui percusserit Philistheum hunc et tulerit obprobrium de Israhel quis est enim hic Philistheus incircumcisus qui expro-

Liber I Samuelis

brauit acies Dei uiuentis 17:27 referebat autem ei populus 17:27 And the people answered him after this manner, say-

ing, So shall it be done to the man that killeth him.

eundem sermonem dicens haec dabuntur uiro qui percusserit eum 17:28 quod cum audisset Heliab frater eius maior 17:28 And Eliab his eldest brother heard when he spake

unto the men; and Eliab’s anger was kindled against David, and he said, Why camest thou down hither? and with whom hast thou left those few sheep in the wilderness? I know thy pride, and the naughtiness of thine heart; for thou art come down that thou mightest see the battle.

loquente eo cum aliis iratus est contra Dauid et ait quare uenisti et quare dereliquisti pauculas oues illas in deserto ego noui superbiam tuam et nequitiam cordis tui quia ut uideres proelium descendisti 17:29 et dixit Dauid quid feci num- 17:29 And David said, What have I now done? Is there

not a cause?

quid non uerbum est 17:30 et declinauit paululum ab eo ad 17:30 And he turned from him toward another, and spake

after the same manner: and the people answered him again after the former manner.

alium dixitque eundem sermonem et respondit ei populus uerbum sicut et prius 17:31 audita sunt autem uerba quae 17:31 And when the words were heard which David spake,

they rehearsed them before Saul: and he sent for him.

locutus est Dauid et adnuntiata in conspectu Saul 17:32 ad 17:32 And David said to Saul, Let no man’s heart fail be-

cause of him; thy servant will go and fight with this Philistine.

quem cum fuisset adductus locutus est ei non concidat cor cuiusquam in eo ego seruus tuus uadam et pugnabo aduersus Philistheum 17:33 et ait Saul ad Dauid non uales resistere 17:33 And Saul said to David, Thou art not able to go

against this Philistine to fight with him: for thou art but a youth, and he a man of war from his youth.

Philistheo isti nec pugnare aduersum eum quia puer es hic autem uir bellator ab adulescentia sua 17:34 dixitque Da- 17:34 And David said unto Saul, Thy servant kept his fa-

ther’s sheep, and there came a lion, and a bear, and took a lamb out of the flock:

uid ad Saul pascebat seruus tuus patris sui gregem et ueniebat leo uel ursus tollebatque arietem de medio gregis 17:35 17:35 And I went out after him, and smote him, and deliv-

ered it out of his mouth: and when he arose against me, I caught him by his beard, and smote him, and slew him.

et sequebar eos et percutiebam eruebamque de ore eorum et illi consurgebant aduersum me et adprehendebam mentum eorum et suffocabam interficiebamque eos 17:36 nam et le- 17:36 Thy servant slew both the lion and the bear: and this

uncircumcised Philistine shall be as one of them, seeing he hath defied the armies of the living God.

onem et ursum interfeci ego seruus tuus erit igitur et Philistheus hic incircumcisus quasi unus ex eis quia ausus est maledicere exercitum Dei uiuentis 17:37 et ait Dauid Do- 17:37 David said moreover, The LORD that delivered me

out of the paw of the lion, and out of the paw of the bear, he will deliver me out of the hand of this Philistine. And

minus qui eruit me de manu leonis et de manu ursi ipse libe- Saul said unto David, Go, and the LORD be with thee. rabit me de manu Philisthei huius dixit autem Saul ad Dauid uade et Dominus tecum sit 17:38 et induit Saul Dauid ues- 17:38 And Saul armed David with his armour, and he put

an helmet of brass upon his head; also he armed him with a coat of mail.

timentis suis et inposuit galeam aeream super caput eius et uestiuit eum lorica 17:39 accinctus ergo Dauid gladio eius 17:39 And David girded his sword upon his armour, and

he assayed to go; for he had not proved it. And David said unto Saul, I cannot go with these; for I have not proved

super ueste sua coepit temptare si armatus posset incedere them. And David put them off him. non enim habebat consuetudinem dixitque Dauid ad Saul non possum sic incedere quia nec usum habeo et deposuit ea 17:40 et tulit baculum suum quem semper habebat in manibus 17:40 And he took his staff in his hand, and chose him

five smooth stones out of the brook, and put them in a shepherd’s bag which he had, even in a scrip; and his sling was in his hand: and he drew near to the Philistine.

Biblia Sacra Vulgata

et elegit sibi quinque limpidissimos lapides de torrente et misit eos in peram pastoralem quam habebat secum et fundam

17:41 And the Philistine came on and drew near unto manu tulit et processit aduersum Philistheum 17:41 ibat au-

David; and the man that bare the shield went before him.

tem Philistheus incedens et adpropinquans aduersum Dauid

17:42 And when the Philistine looked about, and saw et armiger eius ante eum 17:42 cumque inspexisset Philis-

David, he disdained him: for he was but a youth, and ruddy, and of a fair countenance.

theus et uidisset Dauid despexit eum erat enim adulescens

17:43 And the Philistine said unto David, Am I a dog, that rufus et pulcher aspectu 17:43 et dixit Philistheus ad Dauid

thou comest to me with staves? And the Philistine cursed David by his gods.

numquid ego canis sum quod tu uenis ad me cum baculo et

17:44 And the Philistine said to David, Come to me, and maledixit Philistheus Dauid in diis suis 17:44 dixitque ad

I will give thy flesh unto the fowls of the air, and to the beasts of the field.

Dauid ueni ad me et dabo carnes tuas uolatilibus caeli et be-

17:45 Then said David to the Philistine, Thou comest to stiis terrae 17:45 dixit autem Dauid ad Philistheum tu uenis

me with a sword, and with a spear, and with a shield: but I come to thee in the name of the LORD of hosts, the God

of the armies of Israel, whom thou hast defied. ad me cum gladio et hasta et clypeo ego autem uenio ad te

in nomine Domini exercituum Dei agminum Israhel quibus

17:46 This day will the LORD deliver thee into mine hand; exprobrasti 17:46 hodie et dabit te Dominus in manu mea and I will smite thee, and take thine head from thee; and I will give the carcases of the host of the Philistines this day unto the fowls of the air, and to the wild beasts of the earth; that all the earth may know that there is a God in Israel.

et percutiam te et auferam caput tuum a te et dabo cadauer castrorum Philisthim hodie uolatilibus caeli et bestiis terrae

17:47 And all this assembly shall know that the LORD ut sciat omnis terra quia est Deus in Israhel 17:47 et nouerit

saveth not with sword and spear: for the battle is the LORD’s, and he will give you into our hands.

uniuersa ecclesia haec quia non in gladio nec in hasta saluat Dominus ipsius est enim bellum et tradet uos in manus

17:48 And it came to pass, when the Philistine arose, and nostras 17:48 cum ergo surrexisset Philistheus et ueniret et

came, and drew nigh to meet David, that David hastened, and ran toward the army to meet the Philistine.

adpropinquaret contra Dauid festinauit Dauid et cucurrit ad

17:49 And David put his hand in his bag, and took thence pugnam ex aduerso Philisthei 17:49 et misit manum suam

a stone, and slang it, and smote the Philistine in his forehead, that the stone sunk into his forehead; and he fell

upon his face to the earth. in peram tulitque unum lapidem et funda iecit et percussit

Philistheum in fronte et infixus est lapis in fronte eius et ce-

17:50 So David prevailed over the Philistine with a sling cidit in faciem suam super terram 17:50 praeualuitque Da-

and with a stone, and smote the Philistine, and slew him; but there was no sword in the hand of David.

uid aduersus Philistheum in funda et in lapide percussumque Philistheum interfecit cumque gladium non haberet in manu

17:51 Therefore David ran, and stood upon the Philistine, Dauid 17:51 cucurrit et stetit super Philistheum et tulit gla-

and took his sword, and drew it out of the sheath thereof, and slew him, and cut off his head therewith. And when

the Philistines saw their champion was dead, they fled. dium eius et eduxit de uagina sua et interfecit eum praeci-

ditque caput eius uidentes autem Philisthim quod mortuus

17:52 And the men of Israel and of Judah arose, and esset fortissimus eorum fugerunt 17:52 et consurgentes uiri

shouted, and pursued the Philistines, until thou come to the valley, and to the gates of Ekron. And the wounded of the Philistines fell down by the way to Shaaraim, even unto Gath, and unto Ekron.

Israhel et Iuda uociferati sunt et persecuti Philistheos usque dum uenirent in uallem et usque ad portas Accaron cecide-

Liber I Samuelis

runtque uulnerati de Philisthim in uia Sarim usque ad Geth et usque Accaron 17:53 et reuertentes filii Israhel postquam 17:53 And the children of Israel returned from chasing af-

ter the Philistines, and they spoiled their tents.

persecuti fuerant Philistheos inuaserunt castra eorum 17:54 17:54 And David took the head of the Philistine, and

brought it to Jerusalem; but he put his armour in his tent.

adsumens autem Dauid caput Philisthei adtulit illud in Hierusalem arma uero eius posuit in tabernaculo suo 17:55 eo 17:55 And when Saul saw David go forth against the

Philistine, he said unto Abner, the captain of the host, Abner, whose son is this youth? And Abner said, As thy soul

autem tempore quo uiderat Saul Dauid egredientem contra liveth, O king, I cannot tell. Philistheum ait ad Abner principem militiae de qua stirpe descendit hic adulescens Abner dixitque Abner uiuit anima tua rex si noui 17:56 et ait rex interroga tu cuius filius sit iste 17:56 And the king said, Enquire thou whose son the

stripling is.

puer 17:57 cumque regressus esset Dauid percusso Philist- 17:57 And as David returned from the slaughter of the

Philistine, Abner took him, and brought him before Saul with the head of the Philistine in his hand.

heo tulit eum Abner et introduxit coram Saul caput Philisthei habentem in manu 17:58 et ait ad eum Saul de qua progenie 17:58 And Saul said to him, Whose son art thou, thou

young man? And David answered, I am the son of thy servant Jesse the Bethlehemite.

es o adulescens dixitque Dauid filius serui tui Isai Bethleemitae ego sum

18:1 et factum est cum conplesset loqui ad Saul anima Io- 18:1 And it came to pass, when he had made an end of

speaking unto Saul, that the soul of Jonathan was knit with the soul of David, and Jonathan loved him as his own soul.

nathan conligata est animae Dauid et dilexit eum Ionathan quasi animam suam 18:2 tulitque eum Saul in die illa et 18:2 And Saul took him that day, and would let him go no

more home to his father’s house.

non concessit ei ut reuerteretur in domum patris sui 18:3 18:3 Then Jonathan and David made a covenant, because

he loved him as his own soul.

inierunt autem Ionathan et Dauid foedus diligebat enim eum quasi animam suam 18:4 nam expoliauit se Ionathan tuni- 18:4 And Jonathan stripped himself of the robe that was

upon him, and gave it to David, and his garments, even to his sword, and to his bow, and to his girdle.

cam qua erat uestitus et dedit eam Dauid et reliqua uestimenta sua usque ad gladium et arcum suum et usque ad balteum 18:5 egrediebatur quoque Dauid ad omnia quaecum- 18:5 And David went out whithersoever Saul sent him,

and behaved himself wisely: and Saul set him over the men of war, and he was accepted in the sight of all the

que misisset eum Saul et prudenter se agebat posuitque eum people, and also in the sight of Saul’s servants. Saul super uiros belli et acceptus erat in oculis uniuersi populi maximeque in conspectu famulorum Saul 18:6 porro 18:6 And it came to pass as they came, when David was re-

turned from the slaughter of the Philistine, that the women came out of all cities of Israel, singing and dancing, to meet king Saul, with tabrets, with joy, and with instruments of musick.

cum reuerteretur percusso Philistheo Dauid egressae sunt mulieres de uniuersis urbibus Israhel cantantes chorosque ducentes in occursum Saul regis in tympanis laetitiae et in sistris 18:7 et praecinebant mulieres ludentes atque dicen- 18:7 And the women answered one another as they played,

and said, Saul hath slain his thousands, and David his ten thousands.

tes percussit Saul mille et Dauid decem milia 18:8 iratus est

18:8 And Saul was very wroth, and the saying displeased him; and he said, They have ascribed unto David ten thousands, and to me they have ascribed but thousands: and what can he have more but the kingdom?

autem Saul nimis et displicuit in oculis eius iste sermo dixitque dederunt Dauid decem milia et mihi dederunt mille quid

Biblia Sacra Vulgata

18:9 And Saul eyed David from that day and forward. ei superest nisi solum regnum 18:9 non rectis ergo ocu-18:10 And it came to pass on the morrow, that the evil lis Saul aspiciebat Dauid ex die illa et deinceps 18:10 post

spirit from God came upon Saul, and he prophesied in the midst of the house: and David played with his hand, as at

other times: and there was a javelin in Saul’s hand. diem autem alteram inuasit spiritus Dei malus Saul et pro-

phetabat in medio domus suae Dauid autem psallebat manu

18:11 And Saul cast the javelin; for he said, I will smite sua sicut per singulos dies tenebatque Saul lanceam 18:11

David even to the wall with it. And David avoided out of his presence twice.

et misit eam putans quod configere posset Dauid cum pa-

18:12 And Saul was afraid of David, because the LORD riete et declinauit Dauid a facie eius secundo 18:12 et timuit

was with him, and was departed from Saul.

Saul Dauid eo quod esset Dominus cum eo et a se recessisset

18:13 Therefore Saul removed him from him, and made 18:13 amouit ergo eum Saul a se et fecit eum tribunum super

him his captain over a thousand; and he went out and came in before the people.

mille uiros et egrediebatur et intrabat in conspectu populi

18:14 And David behaved himself wisely in all his ways; 18:14 in omnibus quoque uiis suis Dauid prudenter agebat et

and the LORD was with him.

18:15 Wherefore when Saul saw that he behaved himself Dominus erat cum eo 18:15 uidit itaque Saul quod prudens

very wisely, he was afraid of him.

18:16 But all Israel and Judah loved David, because he esset nimis et coepit cauere eum 18:16 omnis autem Israhel

went out and came in before them.

et Iuda diligebat Dauid ipse enim egrediebatur et ingredieba-

18:17 And Saul said to David, Behold my elder daughter tur ante eos 18:17 dixit autem Saul ad Dauid ecce filia mea

Merab, her will I give thee to wife: only be thou valiant for me, and fight the LORD’s battles. For Saul said, Let not mine hand be upon him, but let the hand of the Philistines be upon him.

maior Merob ipsam dabo tibi uxorem tantummodo esto uir fortis et proeliare bella Domini Saul autem reputabat dicens non sit manus mea in eo sed sit super illum manus Philisthi-

18:18 And David said unto Saul, Who am I? and what is norum 18:18 ait autem Dauid ad Saul quis ego sum aut quae

my life, or my father’s family in Israel, that I should be son in law to the king?

est uita mea aut cognatio patris mei in Israhel ut fiam ge-

18:19 But it came to pass at the time when Merab Saul’s ner regis 18:19 factum est autem tempus cum deberet dari

daughter should have been given to David, that she was given unto Adriel the Meholathite to wife.

Merob filia Saul Dauid data est Hadrihel Molathitae uxor

18:20 And Michal Saul’s daughter loved David: and they 18:20 dilexit autem Michol filia Saul altera Dauid et nuntia-

told Saul, and the thing pleased him.

18:21 And Saul said, I will give him her, that she may be tum est Saul et placuit ei 18:21 dixitque Saul dabo eam illi

a snare to him, and that the hand of the Philistines may be against him. Wherefore Saul said to David, Thou shalt

this day be my son in law in the one of the twain. ut fiat ei in scandalum et sit super eum manus Philisthinorum

dixit ergo Saul ad Dauid in duabus rebus gener meus eris ho-

18:22 And Saul commanded his servants, saying, Com- die 18:22 et mandauit Saul seruis suis loquimini ad Dauid

mune with David secretly, and say, Behold, the king hath delight in thee, and all his servants love thee: now there-

fore be the king’s son in law. clam me dicentes ecce places regi et omnes serui eius dili-18:23 And Saul’s servants spake those words in the ears gunt te nunc ergo esto gener regis 18:23 et locuti sunt serui

of David. And David said, Seemeth it to you a light thing to be a king’s son in law, seeing that I am a poor man, and

lightly esteemed? Saul in auribus Dauid omnia uerba haec et ait Dauid num

parum uobis uidetur generum esse regis ego autem sum uir

18:24 And the servants of Saul told him, saying, On this pauper et tenuis 18:24 et renuntiauerunt serui Saul dicen-

manner spake David.

18:25 And Saul said, Thus shall ye say to David, The king tes huiuscemodi uerba locutus est Dauid 18:25 dixit autem

desireth not any dowry, but an hundred foreskins of the Philistines, to be avenged of the king’s enemies. But Saul thought to make David fall by the hand of the Philistines.

Liber I Samuelis

Saul sic loquimini ad Dauid non habet necesse rex sponsalia nisi tantum centum praeputia Philisthinorum ut fiat ultio de inimicis regis porro Saul cogitabat tradere Dauid in manibus Philisthinorum 18:26 cumque renuntiassent serui eius 18:26 And when his servants told David these words, it

pleased David well to be the king’s son in law: and the days were not expired.

Dauid uerba quae diximus placuit sermo in oculis Dauid ut fieret gener regis 18:27 et post dies paucos surgens Dauid 18:27 Wherefore David arose and went, he and his men,

and slew of the Philistines two hundred men; and David brought their foreskins, and they gave them in full tale to the king, that he might be the king’s son in law. And Saul gave him Michal his daughter to wife.

abiit cum uiris qui sub eo erant et percussis Philisthim ducentis uiris adtulit praeputia eorum et adnumerauit ea regi ut esset gener eius dedit itaque ei Saul Michol filiam suam uxorem 18:28 et uidit Saul et intellexit quia Dominus esset 18:28 And Saul saw and knew that the LORD was with

David, and that Michal Saul’s daughter loved him.

cum Dauid Michol autem filia Saul diligebat eum 18:29 et 18:29 And Saul was yet the more afraid of David; and Saul

became David’s enemy continually.

Saul magis coepit timere Dauid factusque est Saul inimicus Dauid cunctis diebus 18:30 et egressi sunt principes Philis- 18:30 Then the princes of the Philistines went forth: and

it came to pass, after they went forth, that David behaved himself more wisely than all the servants of Saul; so that

thinorum a principio autem egressionis eorum prudentius se his name was much set by. gerebat Dauid quam omnes serui Saul et celebre factum est nomen eius nimis

19:1 locutus est autem Saul ad Ionathan filium suum et ad 19:1 And Saul spake to Jonathan his son, and to all his

servants, that they should kill David.

omnes seruos suos ut occiderent Dauid porro Ionathan filius Saul diligebat Dauid ualde 19:2 et indicauit Ionathan 19:2 But Jonathan Saul’s son delighted much in David:

and Jonathan told David, saying, Saul my father seeketh to kill thee: now therefore, I pray thee, take heed to thyself until the morning, and abide in a secret place, and hide thyself:

Dauid dicens quaerit Saul pater meus occidere te quapropter obserua te quaeso mane et manebis clam et absconderis 19:3 ego autem egrediens stabo iuxta patrem meum in agro 19:3 And I will go out and stand beside my father in the

field where thou art, and I will commune with my father of thee; and what I see, that I will tell thee.

ubicumque fueris et ego loquar de te ad patrem meum et quodcumque uidero nuntiabo tibi 19:4 locutus est ergo Io- 19:4 And Jonathan spake good of David unto Saul his fa-

ther, and said unto him, Let not the king sin against his servant, against David; because he hath not sinned against thee, and because his works have been to thee-ward very good:

nathan de Dauid bona ad Saul patrem suum dixitque ad eum ne pecces rex in seruum tuum Dauid quia non peccauit tibi et

opera eius bona sunt tibi ualde 19:5 et posuit animam suam 19:5 For he did put his life in his hand, and slew the Philistine, and the LORD wrought a great salvation for all Israel: thou sawest it, and didst rejoice: wherefore then wilt thou

in manu sua et percussit Philistheum et fecit Dominus salu- sin against innocent blood, to slay David without a cause? tem magnam uniuerso Israhel uidisti et laetatus es quare ergo peccas in sanguine innoxio interficiens Dauid qui est absque culpa 19:6 quod cum audisset Saul placatus uoce Ionathae 19:6 And Saul hearkened unto the voice of Jonathan: and

Saul sware, As the LORD liveth, he shall not be slain.

iurauit uiuit Dominus quia non occidetur 19:7 uocauit ita- 19:7 And Jonathan called David, and Jonathan shewed

him all those things. And Jonathan brought David to Saul, and he was in his presence, as in times past.

que Ionathan Dauid et indicauit ei omnia uerba haec et in-

Biblia Sacra Vulgata

troduxit Ionathan Dauid ad Saul et fuit ante eum sicut fuerat

19:8 And there was war again: and David went out, and heri et nudius tertius 19:8 motum est autem rursus bellum

fought with the Philistines, and slew them with a great slaughter; and they fled from him.

et egressus Dauid pugnauit aduersus Philisthim percussitque

19:9 And the evil spirit from the LORD was upon Saul, as eos plaga magna et fugerunt a facie eius 19:9 et factus est

he sat in his house with his javelin in his hand: and David played with his hand.

spiritus Domini malus in Saul sedebat autem in domo sua et

19:10 And Saul sought to smite David even to the wall tenebat lanceam porro Dauid psallebat in manu sua 19:10

with the javelin: but he slipped away out of Saul’s presence, and he smote the javelin into the wall: and David

fled, and escaped that night. nisusque est Saul configere lancea Dauid in pariete et decli-

nauit Dauid a facie Saul lancea autem casso uulnere perlata

19:11 Saul also sent messengers unto David’s house, to est in parietem et Dauid fugit et saluatus est nocte illa 19:11

watch him, and to slay him in the morning: and Michal David’s wife told him, saying, If thou save not thy life to

night, to morrow thou shalt be slain. misit ergo Saul satellites suos in domum Dauid ut custodirent eum et interficeretur mane quod cum adnuntiasset Dauid Michol uxor sua dicens nisi saluaueris te nocte hac cras

19:12 So Michal let David down through a window: and morieris 19:12 deposuit eum per fenestram porro ille abiit

he went, and fled, and escaped.

19:13 And Michal took an image, and laid it in the bed, et aufugit atque saluatus est 19:13 tulit autem Michol sta-

and put a pillow of goats’ hair for his bolster, and covered it with a cloth.

tuam et posuit eam super lectum et pellem pilosam caprarum

19:14 And when Saul sent messengers to take David, she posuit ad caput eius et operuit eam uestimentis 19:14 misit

said, He is sick.

autem Saul apparitores qui raperent Dauid et responsum est

19:15 And Saul sent the messengers again to see David, quod aegrotaret 19:15 rursumque misit Saul nuntios ut ui-

saying, Bring him up to me in the bed, that I may slay him.

derent Dauid dicens adferte eum ad me in lecto ut occidatur

19:16 And when the messengers were come in, behold, 19:16 cumque uenissent nuntii inuentum est simulacrum su-

there was an image in the bed, with a pillow of goats’ hair for his bolster.

per lectum et pellis caprarum ad caput eius 19:17 dixitque

19:17 And Saul said unto Michal, Why hast thou deceived me so, and sent away mine enemy, that he is escaped? And Michal answered Saul, He said unto me, Let me go; why should I kill thee?

Saul ad Michol quare sic inlusisti mihi et dimisisti inimicum meum ut fugeret et respondit Michol ad Saul quia ipse locu-

19:18 So David fled, and escaped, and came to Samuel to tus est mihi dimitte me alioquin interficiam te 19:18 Dauid

Ramah, and told him all that Saul had done to him. And he and Samuel went and dwelt in Naioth.

autem fugiens saluatus est et uenit ad Samuhel in Ramatha et nuntiauit ei omnia quae fecerat sibi Saul et abierunt ipse

19:19 And it was told Saul, saying, Behold, David is at et Samuhel et morati sunt in Nahioth 19:19 nuntiatum est

Naioth in Ramah.

autem Sauli a dicentibus ecce Dauid in Nahioth in Rama

19:20 And Saul sent messengers to take David: and when 19:20 misit ergo Saul lictores ut raperent Dauid qui cum ui-

they saw the company of the prophets prophesying, and Samuel standing as appointed over them, the Spirit of God was upon the messengers of Saul, and they also prophesied.

dissent cuneum prophetarum uaticinantium et Samuhel stantem super eos factus est etiam in illis spiritus Domini et pro-

19:21 And when it was told Saul, he sent other messen- phetare coeperunt etiam ipsi 19:21 quod cum nuntiatum es-

gers, and they prophesied likewise. And Saul sent messengers again the third time, and they prophesied also.

set Sauli misit alios nuntios prophetauerunt autem et illi et

Liber I Samuelis

rursum Saul misit tertios nuntios qui et ipsi prophetauerunt 19:22 abiit autem etiam ipse in Ramatha et uenit usque ad ci- 19:22 Then went he also to Ramah, and came to a great

well that is in Sechu: and he asked and said, Where are Samuel and David? And one said, Behold, they be at

sternam magnam quae est in Soccho et interrogauit et dixit Naioth in Ramah. in quo loco sunt Samuhel et Dauid dictumque est ei ecce in Nahioth sunt in Rama 19:23 et abiit in Nahioth in Rama et 19:23 And he went thither to Naioth in Ramah: and the

Spirit of God was upon him also, and he went on, and prophesied, until he came to Naioth in Ramah.

factus est etiam super eum spiritus Dei et ambulabat ingrediens et prophetabat usque dum ueniret in Nahioth in Rama 19:24 et expoliauit se etiam ipse uestimentis suis et propheta- 19:24 And he stripped off his clothes also, and prophesied

before Samuel in like manner, and lay down naked all that day and all that night. Wherefore they say, Is Saul also

uit cum ceteris coram Samuhel et cecidit nudus tota die illa among the prophets? et nocte unde et exiuit prouerbium num et Saul inter prophetas

20:1 fugit autem Dauid de Nahioth quae erat in Rama ue- 20:1 And David fled from Naioth in Ramah, and came

and said before Jonathan, What have I done? what is mine iniquity? and what is my sin before thy father, that he

niensque locutus est coram Ionathan quid feci quae est in- seeketh my life? iquitas mea et quod peccatum meum in patrem tuum quia

quaerit animam meam 20:2 qui dixit ei absit non morie- 20:2 And he said unto him, God forbid; thou shalt not die: behold, my father will do nothing either great or small, but that he will shew it me: and why should my father hide this

ris neque enim faciet pater meus quicquam grande uel par- thing from me? it is not so. uum nisi prius indicauerit mihi hunc ergo celauit me pater

meus sermonem tantummodo nequaquam erit istud 20:3 et 20:3 And David sware moreover, and said, Thy father certainly knoweth that I have found grace in thine eyes; and he saith, Let not Jonathan know this, lest he be grieved: but truly as the LORD liveth, and as thy soul liveth, there is but a step between me and death.

iurauit rursum Dauid et ille ait scit profecto pater tuus quia inueni gratiam in oculis tuis et dicet nesciat hoc Ionathan ne forte tristetur quinimmo uiuit Dominus et uiuit anima tua quia uno tantum ut ita dicam gradu ego morsque diui-

dimur 20:4 et ait Ionathan ad Dauid quodcumque dixerit 20:4 Then said Jonathan unto David, Whatsoever thy soul

desireth, I will even do it for thee.

mihi anima tua faciam tibi 20:5 dixit autem Dauid ad Io- 20:5 And David said unto Jonathan, Behold, to morrow is

the new moon, and I should not fail to sit with the king at meat: but let me go, that I may hide myself in the field

nathan ecce kalendae sunt crastino et ego ex more sedere unto the third day at even. soleo iuxta regem ad uescendum dimitte ergo me ut abscondar in agro usque ad uesperam diei tertiae 20:6 si requisierit 20:6 If thy father at all miss me, then say, David earnestly

asked leave of me that he might run to Bethlehem his city: for there is a yearly sacrifice there for all the family.

me pater tuus respondebis ei rogauit me Dauid ut iret celeriter in Bethleem ciuitatem suam quia uictimae sollemnes ibi sunt uniuersis contribulibus eius 20:7 si dixerit bene pax 20:7 If he say thus, It is well; thy servant shall have peace:

but if he be very wroth, then be sure that evil is determined by him.

erit seruo tuo si autem fuerit iratus scito quia conpleta est

malitia eius 20:8 fac ergo misericordiam in seruum tuum 20:8 Therefore thou shalt deal kindly with thy servant; for thou hast brought thy servant into a covenant of the LORD with thee: notwithstanding, if there be in me iniquity, slay

quia foedus Domini me famulum tuum tecum inire fecisti si me thyself; for why shouldest thou bring me to thy father?

Biblia Sacra Vulgata

autem est in me aliqua iniquitas tu me interfice et ad patrem

20:9 And Jonathan said, Far be it from thee: for if I knew tuum ne introducas me 20:9 et ait Ionathan absit hoc a te certainly that evil were determined by my father to come upon thee, then would not I tell it thee?

neque enim fieri potest ut si certo cognouero conpletam pa-

20:10 Then said David to Jonathan, Who shall tell me? or tris mei esse malitiam contra te non adnuntiem tibi 20:10

what if thy father answer thee roughly?

responditque Dauid ad Ionathan quis nuntiabit mihi si quid

20:11 And Jonathan said unto David, Come, and let us go forte responderit tibi pater tuus dure 20:11 et ait Ionathan

out into the field. And they went out both of them into the field.

ad Dauid ueni egrediamur in agrum cumque exissent ambo

20:12 And Jonathan said unto David, O LORD God of in agrum 20:12 ait Ionathan ad Dauid Domine Deus Israhel

Israel, when I have sounded my father about to morrow any time, or the third day, and, behold, if there be good toward David, and I then send not unto thee, and shew it thee;

si inuestigauero sententiam patris mei crastino uel perendie et aliquid boni fuerit super Dauid et non statim misero ad te

20:13 The LORD do so and much more to Jonathan: but if et notum tibi fecero 20:13 haec faciat Dominus Ionathan et

it please my father to do thee evil, then I will shew it thee, and send thee away, that thou mayest go in peace: and the

LORD be with thee, as he hath been with my father. haec augeat si autem perseuerauerit patris mei malitia aduer-

sum te reuelabo aurem tuam et dimittam te ut uadas in pace

20:14 And thou shalt not only while yet I live shew me the et sit Dominus tecum sicut fuit cum patre meo 20:14 et si kindness of the LORD, that I die not:

uixero facies mihi misericordiam Domini si uero mortuus

20:15 But also thou shalt not cut off thy kindness from my fuero 20:15 non auferas misericordiam tuam a domo mea house for ever: no, not when the LORD hath cut off the enemies of David every one from the face of the earth.

usque in sempiternum quando eradicauerit Dominus inimi-

20:16 So Jonathan made a covenant with the house of cos Dauid unumquemque de terra 20:16 pepigit ergo foedus

David, saying, Let the LORD even require it at the hand of David’s enemies.

Ionathan cum domo Dauid et requisiuit Dominus de manu

20:17 And Jonathan caused David to swear again, because inimicorum Dauid 20:17 et addidit Ionathan deierare Dauid

he loved him: for he loved him as he loved his own soul.

eo quod diligeret illum sicut animam enim suam ita diligebat

20:18 Then Jonathan said to David, To morrow is the new eum 20:18 dixitque ad eum Ionathan cras kalendae sunt et

moon: and thou shalt be missed, because thy seat will be empty.

requireris 20:19 requiretur enim sessio tua usque perendie

20:19 And when thou hast stayed three days, then thou shalt go down quickly, and come to the place where thou didst hide thyself when the business was in hand, and shalt remain by the stone Ezel.

descendes ergo festinus et uenies in locum ubi celandus es in die qua operari licet et sedebis iuxta lapidem cui est nomen

20:20 And I will shoot three arrows on the side thereof, as Ezel 20:20 et ego tres sagittas mittam iuxta eum et iaciam though I shot at a mark. 20:21 And, behold, I will send a lad, saying, Go, find out quasi exercens me ad signum 20:21 mittam quoque et puethe arrows. If I expressly say unto the lad, Behold, the arrows are on this side of thee, take them; then come thou: for there is peace to thee, and no hurt; as the LORD liveth. rum dicens ei uade et adfer mihi sagittas 20:22 si dixero 20:22 But if I say thus unto the young man, Behold, the arrows are beyond thee; go thy way: for the LORD hath sent thee away.

puero ecce sagittae intra te sunt tolle eas tu ueni ad me quia pax tibi est et nihil est mali uiuit Dominus si autem sic locutus fuero puero ecce sagittae ultra te sunt uade quia di-

20:23 And as touching the matter which thou and I have misit te Dominus 20:23 de uerbo autem quod locuti fuimus

spoken of, behold, the LORD be between thee and me for ever.

ego et tu sit Dominus inter me et te usque in sempiternum

Liber I Samuelis

20:24 absconditus est ergo Dauid in agro et uenerunt kalen- 20:24 So David hid himself in the field: and when the new

moon was come, the king sat him down to eat meat.

dae et sedit rex ad comedendum panem 20:25 cumque se- 20:25 And the king sat upon his seat, as at other times, even upon a seat by the wall: and Jonathan arose, and Abner sat by Saul’s side, and David’s place was empty.

disset rex super cathedram suam secundum consuetudinem quae erat iuxta parietem surrexit Ionathan et sedit Abner ex

latere Saul uacuusque apparuit locus Dauid 20:26 et non 20:26 Nevertheless Saul spake not any thing that day: for he thought, Something hath befallen him, he is not clean; surely he is not clean.

est locutus Saul quicquam in die illa cogitabat enim quod forte euenisset ei ut non esset mundus nec purificatus 20:27 20:27 And it came to pass on the morrow, which was the

second day of the month, that David’s place was empty: and Saul said unto Jonathan his son, Wherefore cometh

cumque inluxisset dies secunda post kalendas rursum uacuus not the son of Jesse to meat, neither yesterday, nor to day? apparuit locus Dauid dixitque Saul ad Ionathan filium suum cur non uenit filius Isai nec heri nec hodie ad uescendum 20:28 et respondit Ionathan Sauli rogauit me obnixe ut iret 20:28 And Jonathan answered Saul, David earnestly asked

leave of me to go to Bethlehem:

in Bethleem 20:29 et ait dimitte me quoniam sacrificium 20:29 And he said, Let me go, I pray thee; for our family

hath a sacrifice in the city; and my brother, he hath commanded me to be there: and now, if I have found favour in thine eyes, let me get away, I pray thee, and see my brethren. Therefore he cometh not unto the king’s table.

sollemne est in ciuitate unus de fratribus meis accersiuit me nunc ergo si inueni gratiam in oculis tuis uadam cito et uidebo fratres meos ob hanc causam non uenit ad mensam regis 20:30 iratus autem Saul aduersus Ionathan dixit ei fili 20:30 Then Saul’s anger was kindled against Jonathan,

and he said unto him, Thou son of the perverse rebellious woman, do not I know that thou hast chosen the son of Jesse to thine own confusion, and unto the confusion of thy mother’s nakedness?

mulieris uirum ultro rapientis numquid ignoro quia diligis filium Isai in confusionem tuam et in confusionem ignominiosae matris tuae 20:31 omnibus enim diebus quibus fi- 20:31 For as long as the son of Jesse liveth upon the

ground, thou shalt not be established, nor thy kingdom. Wherefore now send and fetch him unto me, for he shall

lius Isai uixerit super terram non stabilieris tu neque regnum surely die. tuum itaque iam nunc mitte et adduc eum ad me quia filius mortis est 20:32 respondens autem Ionathan Sauli patri suo 20:32 And Jonathan answered Saul his father, and said

unto him, Wherefore shall he be slain? what hath he done?

ait quare moritur quid fecit 20:33 et arripuit Saul lanceam 20:33 And Saul cast a javelin at him to smite him:

whereby Jonathan knew that it was determined of his father to slay David.

ut percuteret eum et intellexit Ionathan quod definitum esset patri suo ut interficeret Dauid 20:34 surrexit ergo Ionat- 20:34 So Jonathan arose from the table in fierce anger, and

did eat no meat the second day of the month: for he was grieved for David, because his father had done him shame.

han a mensa in ira furoris et non comedit in die kalendarum secunda panem contristatus est enim super Dauid eo quod confudisset eum pater suus 20:35 cumque inluxisset mane 20:35 And it came to pass in the morning, that Jonathan

went out into the field at the time appointed with David, and a little lad with him.

uenit Ionathan in agrum iuxta placitum Dauid et puer paruulus cum eo 20:36 et ait ad puerum suum uade et adfer mihi 20:36 And he said unto his lad, Run, find out now the ar-

rows which I shoot. And as the lad ran, he shot an arrow beyond him.

sagittas quas ego iacio cumque puer cucurrisset iecit aliam sagittam trans puerum 20:37 uenit itaque puer ad locum ia- 20:37 And when the lad was come to the place of the arrow

which Jonathan had shot, Jonathan cried after the lad, and said, Is not the arrow beyond thee?

culi quod miserat Ionathan et clamauit Ionathan post tergum

Biblia Sacra Vulgata

20:38 And Jonathan cried after the lad, Make speed, haste, pueri et ait ecce ibi est sagitta porro ultra te 20:38 clamauit-

stay not. And Jonathan’s lad gathered up the arrows, and came to his master.

que Ionathan post tergum pueri festina uelociter ne steteris collegit autem puer Ionathae sagittas et adtulit ad dominum

20:39 But the lad knew not any thing: only Jonathan and suum 20:39 et quid ageretur penitus ignorabat tantummodo

David knew the matter.

20:40 And Jonathan gave his artillery unto his lad, and enim Ionathan et Dauid rem nouerant 20:40 dedit igitur Io-

said unto him, Go, carry them to the city.

nathan arma sua puero et dixit ei uade defer in ciuitatem

20:41 And as soon as the lad was gone, David arose out 20:41 cumque abisset puer surrexit Dauid de loco qui uerge-

of a place toward the south, and fell on his face to the ground, and bowed himself three times: and they kissed one another, and wept one with another, until David exceeded.

bat ad austrum et cadens pronus in terram adorauit tertio et osculantes alterutrum fleuerunt pariter Dauid autem amplius

20:42 And Jonathan said to David, Go in peace, forasmuch 20:42 dixit ergo Ionathan ad Dauid uade in pace quaecumas we have sworn both of us in the name of the LORD, saying, The LORD be between me and thee, and between

my seed and thy seed for ever. que iurauimus ambo in nomine Domini dicentes Dominus

sit inter me et te et inter semen meum et semen tuum usque

20:43 And he arose and departed: and Jonathan went into in sempiternum 20:43 et surrexit et abiit sed et Ionathan in-

the city.

gressus est ciuitatem

21:1 Then came David to Nob to Ahimelech the priest: 21:1 uenit autem Dauid in Nobe ad Ahimelech sacerdo-

and Ahimelech was afraid at the meeting of David, and said unto him, Why art thou alone, and no man with thee?

tem et obstipuit Ahimelech eo quod uenisset Dauid et dixit

21:2 And David said unto Ahimelech the priest, The king ei quare tu solus et nullus est tecum 21:2 et ait Dauid ad hath commanded me a business, and hath said unto me, Let no man know any thing of the business whereabout I send thee, and what I have commanded thee: and I have appointed my servants to such and such a place.

Ahimelech sacerdotem rex praecepit mihi sermonem et dixit nemo sciat rem propter quam a me missus es et cuiusmodi tibi praecepta dederim nam et pueris condixi in illum et illum

21:3 Now therefore what is under thine hand? give me five locum 21:3 nunc igitur si quid habes ad manum uel quinque

loaves of bread in mine hand, or what there is present.

21:4 And the priest answered David, and said, There is panes da mihi aut quicquid inueneris 21:4 et respondens sa-

no common bread under mine hand, but there is hallowed bread; if the young men have kept themselves at least from

women. cerdos Dauid ait ei non habeo panes laicos ad manum sed

tantum panem sanctum si mundi sunt pueri maxime a mulie-

21:5 And David answered the priest, and said unto him, ribus 21:5 et respondit Dauid sacerdoti et dixit ei equidem

Of a truth women have been kept from us about these three days, since I came out, and the vessels of the young men are holy, and the bread is in a manner common, yea, though it were sanctified this day in the vessel.

si de mulieribus agitur continuimus nos ab heri et nudius tertius quando egrediebamur et fuerunt uasa puerorum sancta porro uia haec polluta est sed et ipsa hodie sanctificabitur

21:6 So the priest gave him hallowed bread: for there was in uasis 21:6 dedit ergo ei sacerdos sanctificatum panem no bread there but the shewbread, that was taken from before the LORD, to put hot bread in the day when it was

taken away. neque enim erat ibi panis nisi tantum panes propositionis

qui sublati fuerant a facie Domini ut ponerentur panes calidi

21:7 Now a certain man of the servants of Saul was there 21:7 erat autem ibi uir de seruis Saul in die illa intus in taber-

that day, detained before the LORD; and his name was Doeg, an Edomite, the chiefest of the herdmen that be-

longed to Saul. naculo Domini et nomen eius Doec Idumeus potentissimus

Liber I Samuelis

pastorum Saul 21:8 dixit autem Dauid ad Ahimelech si ha- 21:8 And David said unto Ahimelech, And is there not

here under thine hand spear or sword? for I have neither brought my sword nor my weapons with me, because the

bes hic ad manum hastam aut gladium quia gladium meum et king’s business required haste.

arma mea non tuli mecum sermo enim regis urguebat 21:9 21:9 And the priest said, The sword of Goliath the Philistine, whom thou slewest in the valley of Elah, behold, it is here wrapped in a cloth behind the ephod: if thou wilt take that, take it: for there is no other save that here. And David said, There is none like that; give it me.

et dixit sacerdos gladius Goliath Philisthei quem percussisti in ualle Terebinthi est inuolutus pallio post ephod si istum uis tollere tolle neque enim est alius hic absque eo et ait Dauid non est huic alter similis da mihi eum 21:10 surrexit 21:10 And David arose and fled that day for fear of Saul,

and went to Achish the king of Gath.

itaque Dauid et fugit in die illa a facie Saul et uenit ad Achis regem Geth 21:11 dixeruntque ei serui Achis numquid non 21:11 And the servants of Achish said unto him, Is not

this David the king of the land? did they not sing one to another of him in dances, saying, Saul hath slain his

iste est Dauid rex terrae nonne huic cantabant per choros thousands, and David his ten thousands? dicentes percussit Saul mille et Dauid decem milia 21:12 21:12 And David laid up these words in his heart, and was

sore afraid of Achish the king of Gath.

posuit autem Dauid sermones istos in corde suo et extimuit ualde a facie Achis regis Geth 21:13 et inmutauit os suum 21:13 And he changed his behaviour before them, and

feigned himself mad in their hands, and scrabbled on the doors of the gate, and let his spittle fall down upon his

coram eis et conlabebatur inter manus eorum et inpingebat beard. in ostia portae defluebantque saliuae eius in barbam 21:14 21:14 Then said Achish unto his servants, Lo, ye see the

man is mad: wherefore then have ye brought him to me?

et ait Achis ad seruos suos uidistis hominem insanum quare adduxistis eum ad me 21:15 an desunt nobis furiosi quod 21:15 Have I need of mad men, that ye have brought this

fellow to play the mad man in my presence? shall this fellow come into my house?

introduxistis istum ut fureret me praesente hicine ingredietur domum meam

22:1 abiit ergo inde Dauid et fugit in speluncam Odollam 22:1 David therefore departed thence, and escaped to the cave Adullam: and when his brethren and all his father’s house heard it, they went down thither to him.

quod cum audissent fratres eius et omnis domus patris eius

descenderunt ad eum illuc 22:2 et conuenerunt ad eum om- 22:2 And every one that was in distress, and every one that was in debt, and every one that was discontented, gathered themselves unto him; and he became a captain over them:

nes qui erant in angustia constituti et oppressi aere alieno et and there were with him about four hundred men. amaro animo et factus est eorum princeps fueruntque cum eo quasi quadringenti uiri 22:3 et profectus est Dauid inde 22:3 And David went thence to Mizpeh of Moab: and he

said unto the king of Moab, Let my father and my mother, I pray thee, come forth, and be with you, till I know what

in Maspha quae est Moab et dixit ad regem Moab maneat God will do for me. oro pater meus et mater mea uobiscum donec sciam quid faciat mihi Deus 22:4 et reliquit eos ante faciem regis Moab 22:4 And he brought them before the king of Moab: and

they dwelt with him all the while that David was in the hold.

manseruntque apud eum cunctis diebus quibus Dauid fuit in praesidio 22:5 dixitque Gad propheta ad Dauid noli manere 22:5 And the prophet Gad said unto David, Abide not in

the hold; depart, and get thee into the land of Judah. Then David departed, and came into the forest of Hareth.

in praesidio proficiscere et uade in terram Iuda et profectus Dauid uenit in saltum Hareth 22:6 et audiuit Saul quod ap- 22:6 When Saul heard that David was discovered, and the

men that were with him, (now Saul abode in Gibeah under a tree in Ramah, having his spear in his hand, and all his

paruisset Dauid et uiri qui erant cum eo Saul autem cum servants were standing about him;)

Biblia Sacra Vulgata

maneret in Gabaa et esset in nemore quod est in Rama hastam manu tenens cunctique socii eius circumstarent eum

22:7 Then Saul said unto his servants that stood about him, 22:7 ait ad seruos suos qui adsistebant ei audite filii Iemini

Hear now, ye Benjamites; will the son of Jesse give every one of you fields and vineyards, and make you all captains

of thousands, and captains of hundreds; numquid omnibus uobis dabit filius Isai agros et uineas et 22:8 That all of you have conspired against me, and there uniuersos uos faciet tribunos et centuriones 22:8 quoniam

is none that sheweth me that my son hath made a league with the son of Jesse, and there is none of you that is sorry for me, or sheweth unto me that my son hath stirred up my servant against me, to lie in wait, as at this day?

coniurastis omnes aduersum me et non est qui mihi renuntiet maxime cum et filius meus foedus iunxerit cum filio Isai non est qui uicem meam doleat ex uobis nec qui adnuntiet mihi eo quod suscitauerit filius meus seruum meum aduer-

22:9 Then answered Doeg the Edomite, which was set sum me insidiantem mihi usque hodie 22:9 respondens au-

over the servants of Saul, and said, I saw the son of Jesse coming to Nob, to Ahimelech the son of Ahitub.

tem Doec Idumeus qui adsistebat et erat primus inter seruos Saul uidi inquit filium Isai in Nobe apud Ahimelech filium

22:10 And he enquired of the LORD for him, and gave Achitob 22:10 qui consuluit pro eo Dominum et cibaria de-

him victuals, and gave him the sword of Goliath the Philistine.

dit ei sed et gladium Goliath Philisthei dedit illi 22:11 misit

22:11 Then the king sent to call Ahimelech the priest, the son of Ahitub, and all his father’s house, the priests that were in Nob: and they came all of them to the king.

ergo rex ad accersiendum Ahimelech filium Achitob sacerdotem et omnem domum patris eius sacerdotum qui erant

22:12 And Saul said, Hear now, thou son of Ahitub. And in Nobe qui uenerunt uniuersi ad regem 22:12 et ait Saul

he answered, Here I am, my lord.

22:13 And Saul said unto him, Why have ye conspired audi fili Achitob qui respondit praesto sum domine 22:13

against me, thou and the son of Jesse, in that thou hast given him bread, and a sword, and hast enquired of God for him, that he should rise against me, to lie in wait, as at this day?

dixitque ad eum Saul quare coniurastis aduersum me tu et filius Isai et dedisti ei panes et gladium et consuluisti pro eo Deum ut consurgeret aduersum me insidiator usque ho-

22:14 Then Ahimelech answered the king, and said, And die permanens 22:14 respondensque Ahimelech regi ait et

who is so faithful among all thy servants as David, which is the king’s son in law, and goeth at thy bidding, and is

honourable in thine house? quis in omnibus seruis tuis sicut Dauid fidelis et gener re-

gis et pergens ad imperium tuum et gloriosus in domo tua

22:15 Did I then begin to enquire of God for him? be it 22:15 num hodie coepi consulere pro eo Deum absit hoc a

far from me: let not the king impute any thing unto his servant, nor to all the house of my father: for thy servant

knew nothing of all this, less or more. me ne suspicetur rex aduersus seruum suum rem huiuscemodi in uniuersa domo patris mei non enim sciuit seruus tuus quicquam super hoc negotio uel modicum uel grande

22:16 And the king said, Thou shalt surely die, Ahim- 22:16 dixitque rex morte morieris Ahimelech tu et omnis do-

elech, thou, and all thy father’s house.

22:17 And the king said unto the footmen that stood about mus patris tui 22:17 et ait rex emissariis qui circumstabant

him, Turn, and slay the priests of the LORD: because their hand also is with David, and because they knew when he fled, and did not shew it to me. But the servants of the king would not put forth their hand to fall upon the priests of the LORD.

eum conuertimini et interficite sacerdotes Domini nam manus eorum cum Dauid est scientes quod fugisset non indicauerunt mihi noluerunt autem serui regis extendere manum

Liber I Samuelis

suam in sacerdotes Domini 22:18 et ait rex Doec conuer- 22:18 And the king said to Doeg, Turn thou, and fall upon

the priests. And Doeg the Edomite turned, and he fell upon the priests, and slew on that day fourscore and five persons

tere tu et inrue in sacerdotes conuersusque Doec Idumeus that did wear a linen ephod. inruit in sacerdotes et trucidauit in die illa octoginta quinque uiros uestitos ephod lineo 22:19 Nobe autem ciuitatem sa- 22:19 And Nob, the city of the priests, smote he with the

edge of the sword, both men and women, children and sucklings, and oxen, and asses, and sheep, with the edge

cerdotum percussit in ore gladii uiros et mulieres paruulos of the sword. et lactantes bouem et asinum et ouem in ore gladii 22:20 22:20 And one of the sons of Ahimelech the son of Ahitub,

named Abiathar, escaped, and fled after David.

euadens autem unus filius Ahimelech filii Achitob cuius nomen erat Abiathar fugit ad Dauid 22:21 et adnuntiauit ei 22:21 And Abiathar shewed David that Saul had slain the

LORD’s priests.

quod occidisset Saul sacerdotes Domini 22:22 et ait Dauid 22:22 And David said unto Abiathar, I knew it that day,

when Doeg the Edomite was there, that he would surely tell Saul: I have occasioned the death of all the persons of

ad Abiathar sciebam in die illa quod cum ibi esset Doec Idu- thy father’s house. meus procul dubio adnuntiaret Saul ego sum reus omnium animarum patris tui 22:23 mane mecum ne timeas si quis 22:23 Abide thou with me, fear not: for he that seeketh

my life seeketh thy life: but with me thou shalt be in safeguard.

quaesierit animam meam quaeret et animam tuam mecumque seruaberis

23:1 et nuntiauerunt Dauid dicentes ecce Philisthim obpu- 23:1 Then they told David, saying, Behold, the Philistines

fight against Keilah, and they rob the threshingfloors.

gnant Ceila et diripiunt areas 23:2 consuluit igitur Dauid 23:2 Therefore David enquired of the LORD, saying, Shall

I go and smite these Philistines? And the LORD said unto David, Go, and smite the Philistines, and save Keilah.

Dominum dicens num uadam et percutiam Philistheos istos et ait Dominus ad Dauid uade et percuties Philistheos et saluabis Ceila 23:3 et dixerunt uiri qui erant cum Dauid 23:3 And David’s men said unto him, Behold, we be afraid

here in Judah: how much more then if we come to Keilah against the armies of the Philistines?

ad eum ecce nos hic in Iudaea consistentes timemus quanto magis si ierimus in Ceila aduersum agmina Philisthinorum 23:4 rursum ergo Dauid consuluit Dominum qui respondens 23:4 Then David enquired of the LORD yet again. And the

LORD answered him and said, Arise, go down to Keilah; for I will deliver the Philistines into thine hand.

ei ait surge et uade in Ceila ego enim tradam Philistheos in manu tua 23:5 abiit Dauid et uiri eius in Ceila et pugnauit 23:5 So David and his men went to Keilah, and fought

with the Philistines, and brought away their cattle, and smote them with a great slaughter. So David saved the

aduersum Philistheos et abegit iumenta eorum et percussit inhabitants of Keilah. eos plaga magna et saluauit Dauid habitatores Ceilae 23:6 23:6 And it came to pass, when Abiathar the son of Ahim-

elech fled to David to Keilah, that he came down with an ephod in his hand.

porro eo tempore quo fugiebat Abiathar filius Ahimelech ad Dauid in Ceila ephod secum habens descenderat 23:7 nun- 23:7 And it was told Saul that David was come to Keilah.

And Saul said, God hath delivered him into mine hand; for he is shut in, by entering into a town that hath gates and

tiatum est autem Saul quod uenisset Dauid in Ceila et ait bars. Saul tradidit eum Deus in manus meas conclususque est introgressus urbem in qua portae et serae 23:8 et praecepit 23:8 And Saul called all the people together to war, to go

down to Keilah, to besiege David and his men.

Saul omni populo ut ad pugnam descenderet in Ceila et obsideret Dauid et uiros eius 23:9 quod cum rescisset Dauid 23:9 And David knew that Saul secretly practised mis-

chief against him; and he said to Abiathar the priest, Bring hither the ephod.

Biblia Sacra Vulgata

quia praepararet ei Saul clam malum dixit ad Abiathar sacer-

23:10 Then said David, O LORD God of Israel, thy ser- dotem adplica ephod 23:10 et ait Dauid Domine Deus Isra-

vant hath certainly heard that Saul seeketh to come to Keilah, to destroy the city for my sake.

hel audiuit famam seruus tuus quod disponat Saul uenire ad

23:11 Will the men of Keilah deliver me up into his hand? Ceila ut euertat urbem propter me 23:11 si tradent me uiri

will Saul come down, as thy servant hath heard? O LORD God of Israel, I beseech thee, tell thy servant. And the

LORD said, He will come down. Ceila in manus eius et si descendet Saul sicut audiuit seruus

tuus Domine Deus Israhel indica seruo tuo et ait Dominus

23:12 Then said David, Will the men of Keilah deliver me descendet 23:12 dixitque Dauid si tradent uiri Ceilae me et

and my men into the hand of Saul? And the LORD said, They will deliver thee up.

uiros qui sunt mecum in manu Saul et dixit Dominus tradent

23:13 Then David and his men, which were about six hun- 23:13 surrexit ergo Dauid et uiri eius quasi sescenti et egressi

dred, arose and departed out of Keilah, and went whithersoever they could go. And it was told Saul that David was

escaped from Keilah; and he forbare to go forth. de Ceila huc atque illuc uagabantur incerti nuntiatumque est

Saul quod fugisset Dauid de Ceila quam ob rem dissimu-

23:14 And David abode in the wilderness in strong holds, lauit exire 23:14 morabatur autem Dauid in deserto in lo-

and remained in a mountain in the wilderness of Ziph. And Saul sought him every day, but God delivered him

not into his hand. cis firmissimis mansitque in monte solitudinis Ziph quaere-

bat tamen eum Saul cunctis diebus et non tradidit eum De-

23:15 And David saw that Saul was come out to seek his us in manus eius 23:15 et uidit Dauid quod egressus esset

life: and David was in the wilderness of Ziph in a wood.

Saul ut quaereret animam eius porro Dauid erat in deserto

23:16 And Jonathan Saul’s son arose, and went to David Ziph in silua 23:16 et surrexit Ionathan filius Saul et abiit

into the wood, and strengthened his hand in God.

ad Dauid in silua et confortauit manus eius in Deo dixitque

23:17 And he said unto him, Fear not: for the hand of Saul ei 23:17 ne timeas neque enim inueniet te manus Saul pat-

my father shall not find thee; and thou shalt be king over Israel, and I shall be next unto thee; and that also Saul my

father knoweth. ris mei et tu regnabis super Israhel et ego ero tibi secundus 23:18 And they two made a covenant before the LORD: sed et Saul pater meus scit hoc 23:18 percussit igitur uter-

and David abode in the wood, and Jonathan went to his house.

que foedus coram Domino mansitque Dauid in silua Ionat-

23:19 Then came up the Ziphites to Saul to Gibeah, say- has autem reuersus est in domum suam 23:19 ascenderunt

ing, Doth not David hide himself with us in strong holds in the wood, in the hill of Hachilah, which is on the south

of Jeshimon? autem Ziphei ad Saul in Gabaa dicentes nonne Dauid lati-

tat apud nos in locis tutissimis siluae in colle Achilae quae

23:20 Now therefore, O king, come down according to all est ad dexteram deserti 23:20 nunc ergo sicut desiderauit

the desire of thy soul to come down; and our part shall be to deliver him into the king’s hand.

anima tua ut descenderes descende nostrum autem erit ut

23:21 And Saul said, Blessed be ye of the LORD; for ye tradamus eum in manus regis 23:21 dixitque Saul benedicti

have compassion on me.

23:22 Go, I pray you, prepare yet, and know and see his uos a Domino quia doluistis uicem meam 23:22 abite oro et

place where his haunt is, and who hath seen him there: for it is told me that he dealeth very subtilly.

diligentius praeparate et curiosius agite et considerate locum ubi sit pes eius uel quis uiderit eum ibi recogitat enim de me

23:23 See therefore, and take knowledge of all the lurking quod callide insidier ei 23:23 considerate et uidete omnia

places where he hideth himself, and come ye again to me with the certainty, and I will go with you: and it shall come to pass, if he be in the land, that I will search him out throughout all the thousands of Judah.

latibula eius in quibus absconditur et reuertimini ad me ad

Liber I Samuelis

rem certam ut uadam uobiscum quod si etiam in terra se abstruserit perscrutabor eum in cunctis milibus Iuda 23:24 at 23:24 And they arose, and went to Ziph before Saul: but

David and his men were in the wilderness of Maon, in the plain on the south of Jeshimon.

illi surgentes abierunt in Ziph ante Saul Dauid autem et uiri eius erant in deserto Maon in campestribus ad dextram Iesimuth 23:25 iuit ergo Saul et socii eius ad quaerendum et 23:25 Saul also and his men went to seek him. And they

told David; wherefore he came down into a rock, and abode in the wilderness of Maon. And when Saul heard

nuntiatum est Dauid statimque descendit ad petram et uer- that, he pursued after David in the wilderness of Maon. sabatur in deserto Maon quod cum audisset Saul persecutus est Dauid in deserto Maon 23:26 et ibat Saul ad latus montis 23:26 And Saul went on this side of the mountain, and

David and his men on that side of the mountain: and David made haste to get away for fear of Saul; for Saul and his men compassed David and his men round about to take them.

ex parte una Dauid autem et uiri eius erant in latere montis ex parte altera porro Dauid desperabat se posse euadere a facie Saul itaque Saul et uiri eius in modum coronae cingebant Dauid et uiros eius ut caperent eos 23:27 et nuntius 23:27 But there came a messenger unto Saul, saying,

Haste thee, and come; for the Philistines have invaded the land.

uenit ad Saul dicens festina et ueni quoniam infuderunt se Philisthim super terram 23:28 reuersus est ergo Saul desis- 23:28 Wherefore Saul returned from pursuing after David,

and went against the Philistines: therefore they called that place Selahammahlekoth.

tens persequi Dauid et perrexit in occursum Philisthinorum propter hoc uocauerunt locum illum petram Diuidentem

24:1 ascendit ergo Dauid inde et habitauit in locis tutis- 24:1 And David went up from thence, and dwelt in strong

holds at Engedi.

simis Engaddi 24:2 cumque reuersus esset Saul postquam 24:2 And it came to pass, when Saul was returned from

following the Philistines, that it was told him, saying, Behold, David is in the wilderness of Engedi.

persecutus est Philistheos nuntiauerunt ei dicentes ecce Dauid in deserto est Engaddi 24:3 adsumens ergo Saul tria mi- 24:3 Then Saul took three thousand chosen men out of all

Israel, and went to seek David and his men upon the rocks of the wild goats.

lia electorum uirorum ex omni Israhel perrexit ad inuestigandum Dauid et uiros eius etiam super abruptissimas petras quae solis hibicibus peruiae sunt 24:4 et uenit ad caulas 24:4 And he came to the sheepcotes by the way, where

was a cave; and Saul went in to cover his feet: and David and his men remained in the sides of the cave.

quoque ouium quae se offerebant uianti eratque ibi spelunca quam ingressus est Saul ut purgaret uentrem porro Dauid et uiri eius in interiori parte speluncae latebant 24:5 et dixe- 24:5 And the men of David said unto him, Behold the day

of which the LORD said unto thee, Behold, I will deliver thine enemy into thine hand, that thou mayest do to him as it shall seem good unto thee. Then David arose, and cut off the skirt of Saul’s robe privily.

runt serui Dauid ad eum ecce dies de qua locutus est Dominus ad te ego tradam tibi inimicum tuum ut facias ei sicut placuerit in oculis tuis surrexit ergo Dauid et praecidit oram clamydis Saul silenter 24:6 post haec percussit cor suum 24:6 And it came to pass afterward, that David’s heart

smote him, because he had cut off Saul’s skirt.

Dauid eo quod abscidisset oram clamydis Saul 24:7 dixit- 24:7 And he said unto his men, The LORD forbid that

I should do this thing unto my master, the LORD’s anointed, to stretch forth mine hand against him, seeing

que ad uiros suos propitius mihi sit Dominus ne faciam hanc he is the anointed of the LORD. rem domino meo christo Domini ut mittam manum meam in

Biblia Sacra Vulgata

24:8 So David stayed his servants with these words, and eum quoniam christus Domini est 24:8 et confregit Dauid

suffered them not to rise against Saul. But Saul rose up out of the cave, and went on his way.

uiros suos sermonibus et non permisit eos ut consurgerent in Saul porro Saul exsurgens de spelunca pergebat coepto iti-

24:9 David also arose afterward, and went out of the cave, nere 24:9 surrexit autem et Dauid post eum et egressus de and cried after Saul, saying, My lord the king. And when Saul looked behind him, David stooped with his face to

the earth, and bowed himself. spelunca clamauit post tergum Saul dicens domine mi rex et

respexit Saul post se et inclinans se Dauid pronus in terram

24:10 And David said to Saul, Wherefore hearest thou adorauit 24:10 dixitque ad Saul quare audis uerba hominum

men’s words, saying, Behold, David seeketh thy hurt? 24:11 Behold, this day thine eyes have seen how that the loquentium Dauid quaerit malum aduersum te 24:11 ecce LORD had delivered thee to day into mine hand in the cave: and some bade me kill thee: but mine eye spared thee; and I said, I will not put forth mine hand against my lord; for he is the LORD’s anointed.

hodie uiderunt oculi tui quod tradiderit te Dominus in manu mea in spelunca et cogitaui ut occiderem te sed pepercit tibi oculus meus dixi enim non extendam manum meam in do-

24:12 Moreover, my father, see, yea, see the skirt of thy mino meo quia christus Domini est 24:12 quin potius parobe in my hand: for in that I cut off the skirt of thy robe, and killed thee not, know thou and see that there is neither evil nor transgression in mine hand, and I have not sinned against thee; yet thou huntest my soul to take it.

ter mi uide et cognosce oram clamydis tuae in manu mea quoniam cum praeciderem summitatem clamydis tuae nolui extendere manum meam in te animaduerte et uide quoniam non est in manu mea malum neque iniquitas neque peccaui in te tu autem insidiaris animae meae ut auferas eam

24:13 The LORD judge between me and thee, and the 24:13 iudicet Dominus inter me et te et ulciscatur me Domi-

LORD avenge me of thee: but mine hand shall not be upon thee.

nus ex te manus autem mea non sit in te 24:14 sicut et in

24:14 As saith the proverb of the ancients, Wickedness proceedeth from the wicked: but mine hand shall not be upon thee.

prouerbio antiquo dicitur ab impiis egredietur impietas ma-

24:15 After whom is the king of Israel come out? after nus ergo mea non sit in te 24:15 quem sequeris rex Israhel

whom dost thou pursue? after a dead dog, after a flea.

quem persequeris canem mortuum sequeris et pulicem unum

24:16 The LORD therefore be judge, and judge between 24:16 sit Dominus iudex et iudicet inter me et te et uideat et

me and thee, and see, and plead my cause, and deliver me out of thine hand.

diiudicet causam meam et eruat me de manu tua 24:17 cum

24:17 And it came to pass, when David had made an end of speaking these words unto Saul, that Saul said, Is this thy voice, my son David? And Saul lifted up his voice, and wept.

autem conplesset Dauid loquens sermones huiuscemodi ad Saul dixit Saul numquid uox haec tua est fili mi Dauid et

24:18 And he said to David, Thou art more righteous than leuauit Saul uocem suam et fleuit 24:18 dixitque ad Dauid

I: for thou hast rewarded me good, whereas I have rewarded thee evil.

iustior tu es quam ego tu enim tribuisti mihi bona ego au-

24:19 And thou hast shewed this day how that thou hast tem reddidi tibi mala 24:19 et tu indicasti hodie quae feceris

dealt well with me: forasmuch as when the LORD had delivered me into thine hand, thou killedst me not.

mihi bona quomodo tradiderit me Dominus in manu tua et

24:20 For if a man find his enemy, will he let him go well non occideris me 24:20 quis enim cum inuenerit inimicum

away? wherefore the LORD reward thee good for that thou hast done unto me this day.

suum dimittet eum in uia bona sed Dominus reddat tibi uicis-

24:21 And now, behold, I know well that thou shalt surely situdinem hanc pro eo quod hodie operatus es in me 24:21 be king, and that the kingdom of Israel shall be established in thine hand.

Liber I Samuelis

et nunc quia scio quod certissime regnaturus sis et habiturus in manu tua regnum Israhel 24:22 iura mihi in Domino ne 24:22 Swear now therefore unto me by the LORD, that

thou wilt not cut off my seed after me, and that thou wilt not destroy my name out of my father’s house.

deleas semen meum post me neque auferas nomen meum de domo patris mei 24:23 et iurauit Dauid Sauli abiit ergo Saul 24:23 And David sware unto Saul. And Saul went home;

but David and his men gat them up unto the hold.

in domum suam et Dauid et uiri eius ascenderunt ad tutiora loca

25:1 mortuus est autem Samuhel et congregatus est uni- 25:1 And Samuel died; and all the Israelites were gathered together, and lamented him, and buried him in his house at Ramah. And David arose, and went down to the wilder-

uersus Israhel et planxerunt eum et sepelierunt in domo sua ness of Paran. in Rama consurgensque Dauid descendit in desertum Pharan 25:2 erat autem uir quispiam in solitudine Maon et possessio 25:2 And there was a man in Maon, whose possessions

were in Carmel; and the man was very great, and he had three thousand sheep, and a thousand goats: and he was

eius in Carmelo et homo ille magnus nimis erantque ei oues shearing his sheep in Carmel. tria milia et mille caprae et accidit ut tonderetur grex eius in Carmelo 25:3 nomen autem uiri illius erat Nabal et no- 25:3 Now the name of the man was Nabal; and the name

of his wife Abigail: and she was a woman of good understanding, and of a beautiful countenance: but the man was churlish and evil in his doings; and he was of the house of Caleb.

men uxoris eius Abigail eratque mulier illa prudentissima et speciosa porro uir eius durus et pessimus et malitiosus erat autem de genere Chaleb 25:4 cum ergo audisset Dauid in 25:4 And David heard in the wilderness that Nabal did

shear his sheep.

deserto quod tonderet Nabal gregem suum 25:5 misit de- 25:5 And David sent out ten young men, and David said

unto the young men, Get you up to Carmel, and go to Nabal, and greet him in my name:

cem iuuenes et dixit eis ascendite in Carmelum et uenietis

ad Nabal et salutabitis eum ex nomine meo pacifice 25:6 et 25:6 And thus shall ye say to him that liveth in prosperity, Peace be both to thee, and peace be to thine house, and peace be unto all that thou hast.

dicetis sic fratribus meis et tibi pax et domui tuae pax et omnibus quaecumque habes sit pax 25:7 audiui quod tonde- 25:7 And now I have heard that thou hast shearers: now

thy shepherds which were with us, we hurt them not, neither was there ought missing unto them, all the while they

rent pastores tui qui erant nobiscum in deserto numquam eis were in Carmel. molesti fuimus nec aliquando defuit eis quicquam de grege omni tempore quo fuerunt nobiscum in Carmelo 25:8 in- 25:8 Ask thy young men, and they will shew thee. Where-

fore let the young men find favour in thine eyes: for we come in a good day: give, I pray thee, whatsoever cometh

terroga pueros tuos et indicabunt tibi nunc ergo inueniant to thine hand unto thy servants, and to thy son David. pueri gratiam in oculis tuis in die enim bona uenimus quodcumque inuenerit manus tua da seruis tuis et filio tuo Dauid 25:9 cumque uenissent pueri Dauid locuti sunt ad Na- 25:9 And when David’s young men came, they spake to

Nabal according to all those words in the name of David, and ceased.

bal omnia uerba haec ex nomine Dauid et siluerunt 25:10

25:10 And Nabal answered David’s servants, and said, Who is David? and who is the son of Jesse? there be many servants now a days that break away every man from his master.

respondens autem Nabal pueris Dauid ait quis est Dauid et quis est filius Isai hodie increuerunt serui qui fugiunt dominos suos 25:11 tollam ergo panes meos et aquas meas 25:11 Shall I then take my bread, and my water, and my

flesh that I have killed for my shearers, and give it unto men, whom I know not whence they be?

et carnes pecorum quae occidi tonsoribus meis et dabo ui-

Biblia Sacra Vulgata

25:12 So David’s young men turned their way, and went ris quos nescio unde sint 25:12 regressi sunt itaque pueri

again, and came and told him all those sayings.

Dauid per uiam suam et reuersi uenerunt et nuntiauerunt ei

25:13 And David said unto his men, Gird ye on every man omnia uerba quae dixerat 25:13 tunc Dauid ait uiris suis ac-

his sword. And they girded on every man his sword; and David also girded on his sword: and there went up after David about four hundred men; and two hundred abode by the stuff.

cingatur unusquisque gladio suo et accincti sunt singuli gladio suo accinctusque est et Dauid ense suo et secuti sunt Dauid quasi quadringenti uiri porro ducenti remanserunt ad

25:14 But one of the young men told Abigail, Nabal’s sarcinas 25:14 Abigail autem uxori Nabal nuntiauit unus de

wife, saying, Behold, David sent messengers out of the wilderness to salute our master; and he railed on them.

pueris dicens ecce misit Dauid nuntios de deserto ut benedi-

25:15 But the men were very good unto us, and we were cerent domino nostro et auersus est eos 25:15 homines isti

not hurt, neither missed we any thing, as long as we were conversant with them, when we were in the fields:

boni satis fuerunt nobis et non molesti nec quicquam aliquando periit omni tempore quo sumus conuersati cum eis

25:16 They were a wall unto us both by night and day, all in deserto 25:16 pro muro erant nobis tam in nocte quam in

the while we were with them keeping the sheep.

25:17 Now therefore know and consider what thou wilt die omnibus diebus quibus pauimus apud eos greges 25:17

do; for evil is determined against our master, and against all his household: for he is such a son of Belial, that a man

cannot speak to him. quam ob rem considera et recogita quid facias quoniam con-

pleta est malitia aduersum uirum tuum et aduersus domum

25:18 Then Abigail made haste, and took two hundred tuam et ipse filius est Belial ita ut nemo ei possit loqui 25:18

loaves, and two bottles of wine, and five sheep ready dressed, and five measures of parched corn, and an hundred clusters of raisins, and two hundred cakes of figs, and laid them on asses.

festinauit igitur Abigail et tulit ducentos panes et duos utres uini et quinque arietes coctos et quinque sata pulentae et centum ligaturas uuae passae et ducentas massas caricarum

25:19 And she said unto her servants, Go on before me; et inposuit super asinos 25:19 dixitque pueris suis praece-

behold, I come after you. But she told not her husband Nabal.

dite me ecce ego post tergum sequar uos uiro autem suo Na-

25:20 And it was so, as she rode on the ass, that she came bal non indicauit 25:20 cum ergo ascendisset asinum et de-

down by the covert on the hill, and, behold, David and his men came down against her; and she met them.

scenderet ad radices montis Dauid et uiri eius descendebant

25:21 Now David had said, Surely in vain have I kept all in occursum eius quibus et illa occurrit 25:21 et ait Dauid

that this fellow hath in the wilderness, so that nothing was missed of all that pertained unto him: and he hath requited

me evil for good. uere frustra seruaui omnia quae huius erant in deserto et non

periit quicquam de cunctis quae ad eum pertinebant et red-

25:22 So and more also do God unto the enemies of David, didit mihi malum pro bono 25:22 haec faciat Deus inimicis

if I leave of all that pertain to him by the morning light any that pisseth against the wall.

Dauid et haec addat si reliquero de omnibus quae ad eum

25:23 And when Abigail saw David, she hasted, and pertinent usque mane mingentem ad parietem 25:23 cum

lighted off the ass, and fell before David on her face, and bowed herself to the ground,

autem uidisset Abigail Dauid festinauit et descendit de asino et procidit coram Dauid super faciem suam et adorauit super

25:24 And fell at his feet, and said, Upon me, my lord, terram 25:24 et cecidit ad pedes eius et dixit in me sit do-

upon me let this iniquity be: and let thine handmaid, I pray thee, speak in thine audience, and hear the words of

thine handmaid. mine mi haec iniquitas loquatur obsecro ancilla tua in auri-

Liber I Samuelis

bus tuis et audi uerba famulae tuae 25:25 ne ponat oro do- 25:25 Let not my lord, I pray thee, regard this man of Be-

lial, even Nabal: for as his name is, so is he; Nabal is his name, and folly is with him: but I thine handmaid saw not

minus meus rex cor suum super uirum istum iniquum Na- the young men of my lord, whom thou didst send. bal quia secundum nomen suum stultus est et est stultitia cum eo ego autem ancilla tua non uidi pueros tuos domine mi quos misisti 25:26 nunc ergo domine mi uiuit Dominus 25:26 Now therefore, my lord, as the LORD liveth, and

as thy soul liveth, seeing the LORD hath withholden thee from coming to shed blood, and from avenging thyself with thine own hand, now let thine enemies, and they that seek evil to my lord, be as Nabal.

et uiuit anima tua qui prohibuit te ne uenires in sanguine et saluauit manum tuam tibi et nunc fiant sicut Nabal inimici tui et qui quaerunt domino meo malum 25:27 quapro- 25:27 And now this blessing which thine handmaid hath

brought unto my lord, let it even be given unto the young men that follow my lord.

pter suscipe benedictionem hanc quam adtulit ancilla tua tibi domino meo et da pueris qui sequuntur te dominum meum 25:28 aufer iniquitatem famulae tuae faciens enim faciet tibi 25:28 I pray thee, forgive the trespass of thine handmaid:

for the LORD will certainly make my lord a sure house; because my lord fighteth the battles of the LORD, and evil

Dominus domino meo domum fidelem quia proelia Domini hath not been found in thee all thy days. domine mi tu proeliaris malitia ergo non inueniatur in te omnibus diebus uitae tuae 25:29 si enim surrexerit aliquando 25:29 Yet a man is risen to pursue thee, and to seek thy

soul: but the soul of my lord shall be bound in the bundle of life with the LORD thy God; and the souls of thine enemies, them shall he sling out, as out of the middle of a sling.

homo persequens te et quaerens animam tuam erit anima domini mei custodita quasi in fasciculo uiuentium apud Dominum Deum tuum porro anima inimicorum tuorum rotabitur quasi in impetu et circulo fundae 25:30 cum ergo fece- 25:30 And it shall come to pass, when the LORD shall

have done to my lord according to all the good that he hath spoken concerning thee, and shall have appointed

rit tibi Dominus domino meo omnia quae locutus est bona thee ruler over Israel; de te et constituerit te ducem super Israhel 25:31 non erit 25:31 That this shall be no grief unto thee, nor offence of

heart unto my lord, either that thou hast shed blood causeless, or that my lord hath avenged himself: but when the LORD shall have dealt well with my lord, then remember thine handmaid.

tibi hoc in singultum et in scrupulum cordis domino meo quod effuderis sanguinem innoxium aut ipse te ultus fueris et cum benefecerit Dominus domino meo recordaberis ancillae tuae 25:32 et ait Dauid ad Abigail benedictus Do- 25:32 And David said to Abigail, Blessed be the LORD

God of Israel, which sent thee this day to meet me:

minus Deus Israhel qui misit te hodie in occursum meum et benedictum eloquium tuum 25:33 et benedicta tu quae 25:33 And blessed be thy advice, and blessed be thou,

which hast kept me this day from coming to shed blood, and from avenging myself with mine own hand.

prohibuisti me hodie ne irem ad sanguinem et ulciscerer me manu mea 25:34 alioquin uiuit Dominus Deus Israhel qui 25:34 For in very deed, as the LORD God of Israel liveth,

which hath kept me back from hurting thee, except thou hadst hasted and come to meet me, surely there had not been left unto Nabal by the morning light any that pisseth against the wall.

prohibuit me malum facere tibi nisi cito uenisses in occursum mihi non remansisset Nabal usque ad lucem matutinam mingens ad parietem 25:35 suscepit ergo Dauid de manu 25:35 So David received of her hand that which she had

brought him, and said unto her, Go up in peace to thine house; see, I have hearkened to thy voice, and have ac-

eius omnia quae adtulerat ei dixitque ei uade pacifice in do- cepted thy person. mum tuam ecce audiui uocem tuam et honoraui faciem tuam

Biblia Sacra Vulgata

25:36 And Abigail came to Nabal; and, behold, he held a 25:36 uenit autem Abigail ad Nabal et ecce erat ei conuiuium

feast in his house, like the feast of a king; and Nabal’s heart was merry within him, for he was very drunken: wherefore she told him nothing, less or more, until the morning light.

in domo eius quasi conuiuium regis et cor Nabal iucundum erat enim ebrius nimis et non indicauit ei uerbum pusillum

25:37 But it came to pass in the morning, when the wine aut grande usque in mane 25:37 diluculo autem cum diges-

was gone out of Nabal, and his wife had told him these things, that his heart died within him, and he became as a

stone. sisset uinum Nabal indicauit ei uxor sua uerba haec et emor-25:38 And it came to pass about ten days after, that the tuum est cor eius intrinsecus et factus est quasi lapis 25:38

LORD smote Nabal, that he died.

cumque pertransissent decem dies percussit Dominus Na-

25:39 And when David heard that Nabal was dead, he said, bal et mortuus est 25:39 quod cum audisset Dauid mortuum

Blessed be the LORD, that hath pleaded the cause of my reproach from the hand of Nabal, and hath kept his servant from evil: for the LORD hath returned the wickedness of Nabal upon his own head. And David sent and communed with Abigail, to take her to him to wife.

Nabal ait benedictus Dominus qui iudicauit causam obprobrii mei de manu Nabal et seruum suum custodiuit a malo et malitiam Nabal reddidit Dominus in caput eius misit ergo Dauid et locutus est ad Abigail ut sumeret eam sibi in uxo-

25:40 And when the servants of David were come to Abi- rem 25:40 et uenerunt pueri Dauid ad Abigail in Carmelum

gail to Carmel, they spake unto her, saying, David sent us unto thee, to take thee to him to wife.

et locuti sunt ad eam dicentes Dauid misit nos ad te ut acci-

25:41 And she arose, and bowed herself on her face to the piat te sibi in uxorem 25:41 quae consurgens adorauit pro-

earth, and said, Behold, let thine handmaid be a servant to wash the feet of the servants of my lord.

na in terram et ait ecce famula tua sit in ancillam ut lauet

25:42 And Abigail hasted, and arose and rode upon an ass, pedes seruorum domini mei 25:42 et festinauit et surrexit

with five damsels of hers that went after her; and she went after the messengers of David, and became his wife.

Abigail et ascendit super asinum et quinque puellae ierunt cum ea pedisequae eius et secuta est nuntios Dauid et facta

25:43 David also took Ahinoam of Jezreel; and they were est illi uxor 25:43 sed et Ahinoem accepit Dauid de Iezra-

also both of them his wives.

25:44 But Saul had given Michal his daughter, David’s hel et fuit utraque uxor eius 25:44 Saul autem dedit Michol

wife, to Phalti the son of Laish, which was of Gallim.

filiam suam uxorem Dauid Falti filio Lais qui erat de Gallim

26:1 And the Ziphites came unto Saul to Gibeah, saying, 26:1 et uenerunt Ziphei ad Saul in Gabaa dicentes ecce Da-

Doth not David hide himself in the hill of Hachilah, which is before Jeshimon?

uid absconditus est in colle Achilae quae est ex aduerso so-

26:2 Then Saul arose, and went down to the wilderness litudinis 26:2 et surrexit Saul et descendit in desertum Ziph

of Ziph, having three thousand chosen men of Israel with him, to seek David in the wilderness of Ziph.

et cum eo tria milia uirorum de electis Israhel ut quaereret

26:3 And Saul pitched in the hill of Hachilah, which is be- Dauid in deserto Ziph 26:3 et castrametatus est Saul in Ga-

fore Jeshimon, by the way. But David abode in the wilderness, and he saw that Saul came after him into the wilder-

ness. baa Achilae quae erat ex aduerso solitudinis in uia Dauid

autem habitabat in deserto uidens autem quod uenisset Saul

26:4 David therefore sent out spies, and understood that post se in desertum 26:4 misit exploratores et didicit quod

Saul was come in very deed.

26:5 And David arose, and came to the place where Saul uenisset certissime 26:5 et surrexit Dauid et uenit ad locum

had pitched: and David beheld the place where Saul lay, and Abner the son of Ner, the captain of his host: and Saul lay in the trench, and the people pitched round about him. ubi erat Saul cumque uidisset locum in quo dormiebat Saul

et Abner filius Ner princeps militiae eius Saulem dormien-

Liber I Samuelis

tem in tentorio et reliquum uulgus per circuitum eius 26:6 26:6 Then answered David and said to Ahimelech the Hit-

tite, and to Abishai the son of Zeruiah, brother to Joab, saying, Who will go down with me to Saul to the camp?

ait Dauid ad Ahimelech Cettheum et Abisai filium Saruiae And Abishai said, I will go down with thee. fratrem Ioab dicens quis descendet mecum ad Saul in castra dixitque Abisai ego descendam tecum 26:7 uenerunt ergo 26:7 So David and Abishai came to the people by night:

and, behold, Saul lay sleeping within the trench, and his spear stuck in the ground at his bolster: but Abner and the

Dauid et Abisai ad populum nocte et inuenerunt Saul iacen- people lay round about him. tem et dormientem in tentorio et hastam fixam in terra ad caput eius Abner autem et populum dormientes in circuitu eius 26:8 dixitque Abisai ad Dauid conclusit Deus hodie inimi- 26:8 Then said Abishai to David, God hath delivered thine

enemy into thine hand this day: now therefore let me smite him, I pray thee, with the spear even to the earth at once,

cum tuum in manus tuas nunc ergo perfodiam eum lancea in and I will not smite him the second time. terra semel et secundo opus non erit 26:9 et dixit Dauid ad 26:9 And David said to Abishai, Destroy him not: for who

can stretch forth his hand against the LORD’s anointed, and be guiltless?

Abisai ne interficias eum quis enim extendit manum suam in christum Domini et innocens erit 26:10 et dixit Dauid 26:10 David said furthermore, As the LORD liveth, the

LORD shall smite him; or his day shall come to die; or he shall descend into battle, and perish.

uiuit Dominus quia nisi Dominus percusserit eum aut dies eius uenerit ut moriatur aut in proelium descendens perierit 26:11 propitius mihi sit Dominus ne extendam manum 26:11 The LORD forbid that I should stretch forth mine

hand against the LORD’s anointed: but, I pray thee, take thou now the spear that is at his bolster, and the cruse of

meam in christum Domini nunc igitur tolle hastam quae est water, and let us go. ad caput eius et scyphum aquae et abeamus 26:12 tulit ergo 26:12 So David took the spear and the cruse of water from

Saul’s bolster; and they gat them away, and no man saw it, nor knew it, neither awaked: for they were all asleep; be-

Dauid hastam et scyphum aquae qui erat ad caput Saul et cause a deep sleep from the LORD was fallen upon them. abierunt et non erat quisquam qui uideret et intellegeret et uigilaret sed omnes dormiebant quia sopor Domini inruerat super eos 26:13 cumque transisset Dauid ex aduerso et ste- 26:13 Then David went over to the other side, and stood

on the top of an hill afar off; a great space being between them:

tisset in uertice montis de longe et esset grande interuallum inter eos 26:14 clamauit Dauid ad populum et ad Abner fi- 26:14 And David cried to the people, and to Abner the son

of Ner, saying, Answerest thou not, Abner? Then Abner answered and said, Who art thou that criest to the king?

lium Ner dicens nonne respondebis Abner et respondens Abner ait quis es tu qui clamas et inquietas regem 26:15 et ait 26:15 And David said to Abner, Art not thou a valiant

man? and who is like to thee in Israel? wherefore then hast thou not kept thy lord the king? for there came one of

Dauid ad Abner numquid non uir tu es et quis alius similis the people in to destroy the king thy lord. tui in Israhel quare ergo non custodisti dominum tuum regem ingressus est enim unus de turba ut interficeret regem dominum tuum 26:16 non est bonum hoc quod fecisti uiuit 26:16 This thing is not good that thou hast done. As the

LORD liveth, ye are worthy to die, because ye have not kept your master, the LORD’s anointed. And now see where the king’s spear is, and the cruse of water that was at his bolster.

Dominus quoniam filii mortis estis uos qui non custodistis dominum uestrum christum Domini nunc ergo uide ubi sit hasta regis et ubi scyphus aquae qui erat ad caput eius 26:17 26:17 And Saul knew David’s voice, and said, Is this thy

voice, my son David? And David said, It is my voice, my lord, O king.

cognouit autem Saul uocem Dauid et dixit num uox tua est

Biblia Sacra Vulgata

26:18 And he said, Wherefore doth my lord thus pursue haec fili mi Dauid et Dauid uox mea domine mi rex 26:18 et

after his servant? for what have I done? or what evil is in mine hand?

ait quam ob causam dominus meus persequitur seruum suum

26:19 Now therefore, I pray thee, let my lord the king hear quid feci aut quod est in manu mea malum 26:19 nunc ergo the words of his servant. If the LORD have stirred thee up against me, let him accept an offering: but if they be the children of men, cursed be they before the LORD; for they have driven me out this day from abiding in the inheritance of the LORD, saying, Go, serve other gods.

audi oro domine mi rex uerba serui tui si Dominus incitat te aduersum me odoretur sacrificium si autem filii hominum maledicti sunt in conspectu Domini qui eiecerunt me hodie ut non habitem in hereditate Domini dicentes uade serui diis

26:20 Now therefore, let not my blood fall to the earth alienis 26:20 et nunc non effundatur sanguis meus in terra

before the face of the LORD: for the king of Israel is come out to seek a flea, as when one doth hunt a partridge in the

mountains. coram Domino quia egressus est rex Israhel ut quaerat puli-26:21 Then said Saul, I have sinned: return, my son David: cem unum sicut persequitur perdix in montibus 26:21 et ait

for I will no more do thee harm, because my soul was precious in thine eyes this day: behold, I have played the

fool, and have erred exceedingly. Saul peccaui reuertere fili mi Dauid nequaquam enim ultra male tibi faciam eo quod pretiosa fuerit anima mea in oculis tuis hodie apparet quod stulte egerim et ignorauerim multa

26:22 And David answered and said, Behold the king’s nimis 26:22 et respondens Dauid ait ecce hasta regis trans-

spear! and let one of the young men come over and fetch it.

eat unus de pueris et tollat eam 26:23 Dominus autem ret-

26:23 The LORD render to every man his righteousness and his faithfulness; for the LORD delivered thee into my hand to day, but I would not stretch forth mine hand against the LORD’s anointed.

ribuet unicuique secundum iustitiam suam et fidem tradidit enim te Dominus hodie in manu mea et nolui leuare manum

26:24 And, behold, as thy life was much set by this day in meam in christum Domini 26:24 et sicuti magnificata est mine eyes, so let my life be much set by in the eyes of the LORD, and let him deliver me out of all tribulation.

anima tua hodie in oculis meis sic magnificetur anima mea

26:25 Then Saul said to David, Blessed be thou, my son in oculis Domini et liberet me de omni angustia 26:25 ait

David: thou shalt both do great things, and also shalt still prevail. So David went on his way, and Saul returned to

his place. ergo Saul ad Dauid benedictus tu fili mi Dauid et quidem faciens facies et potens poteris abiit autem Dauid in uiam suam et Saul reuersus est in locum suum

27:1 And David said in his heart, I shall now perish one 27:1 et ait Dauid in corde suo aliquando incidam in uno die

day by the hand of Saul: there is nothing better for me than that I should speedily escape into the land of the Philistines; and Saul shall despair of me, to seek me any more in any coast of Israel: so shall I escape out of his hand.

in manu Saul nonne melius est ut fugiam et saluer in terra Philisthinorum ut desperet Saul cessetque me quaerere in

27:2 And David arose, and he passed over with the six cunctis finibus Israhel fugiam ergo manus eius 27:2 et sur-

hundred men that were with him unto Achish, the son of Maoch, king of Gath.

rexit Dauid et abiit ipse et sescenti uiri cum eo ad Achis fi-

27:3 And David dwelt with Achish at Gath, he and his lium Mahoc regem Geth 27:3 et habitauit Dauid cum Achis

men, every man with his household, even David with his two wives, Ahinoam the Jezreelitess, and Abigail the

Carmelitess, Nabal’s wife. in Geth ipse et uiri eius uir et domus eius Dauid et duae uxo-

res eius Ahinoem Iezrahelites et Abigail uxor Nabal Carmeli

27:4 And it was told Saul that David was fled to Gath: and 27:4 et nuntiatum est Saul quod fugisset Dauid in Geth et

he sought no more again for him.

27:5 And David said unto Achish, If I have now found non addidit ultra ut quaereret eum 27:5 dixit autem Dauid

grace in thine eyes, let them give me a place in some town in the country, that I may dwell there: for why should thy servant dwell in the royal city with thee?

Liber I Samuelis

ad Achis si inueni gratiam in oculis tuis detur mihi locus in una urbium regionis huius ut habitem ibi cur enim manet seruus tuus in ciuitate regis tecum 27:6 dedit itaque ei 27:6 Then Achish gave him Ziklag that day: wherefore

Ziklag pertaineth unto the kings of Judah unto this day.

Achis in die illa Siceleg propter quam causam facta est Siceleg regum Iuda usque in diem hanc 27:7 fuit autem nu- 27:7 And the time that David dwelt in the country of the

Philistines was a full year and four months.

merus dierum quibus habitauit Dauid in regione Philisthinorum quattuor mensuum 27:8 et ascendit Dauid et uiri eius 27:8 And David and his men went up, and invaded the

Geshurites, and the Gezrites, and the Amalekites: for those nations were of old the inhabitants of the land, as

et agebant praedas de Gesuri et de Gedri et de Amalechitis thou goest to Shur, even unto the land of Egypt. hii enim pagi habitabantur in terra antiquitus euntibus Sur usque ad terram Aegypti 27:9 et percutiebat Dauid omnem 27:9 And David smote the land, and left neither man nor

woman alive, and took away the sheep, and the oxen, and the asses, and the camels, and the apparel, and returned,

terram nec relinquebat uiuentem uirum et mulierem tollens- and came to Achish. que oues et boues et asinos et camelos et uestes reuertebatur et ueniebat ad Achis 27:10 dicebat autem ei Achis in quem 27:10 And Achish said, Whither have ye made a road to

day? And David said, Against the south of Judah, and against the south of the Jerahmeelites, and against the

inruisti hodie respondebatque Dauid contra meridiem Iudae south of the Kenites. et contra meridiem Hiramel et contra meridiem Ceni 27:11 27:11 And David saved neither man nor woman alive, to

bring tidings to Gath, saying, Lest they should tell on us, saying, So did David, and so will be his manner all the

uirum et mulierem non uiuificabat Dauid nec adducebat in while he dwelleth in the country of the Philistines. Geth dicens ne forte loquantur aduersum nos haec fecit Dauid et hoc erat decretum illi omnibus diebus quibus habitauit in regione Philisthinorum 27:12 credidit ergo Achis Dauid 27:12 And Achish believed David, saying, He hath made

his people Israel utterly to abhor him; therefore he shall be my servant for ever.

dicens multa mala operatus est contra populum suum Israhel erit igitur mihi seruus sempiternus

28:1 factum est autem in diebus illis congregauerunt Phi- 28:1 And it came to pass in those days, that the Philistines

gathered their armies together for warfare, to fight with Israel. And Achish said unto David, Know thou assuredly,

listhim agmina sua ut praepararentur ad bellum contra Is- that thou shalt go out with me to battle, thou and thy men. rahel dixitque Achis ad Dauid sciens nunc scito quoniam mecum egredieris in castris tu et uiri tui 28:2 dixitque Da- 28:2 And David said to Achish, Surely thou shalt know

what thy servant can do. And Achish said to David, Therefore will I make thee keeper of mine head for ever.

uid ad Achis nunc scies quae facturus est seruus tuus et ait Achis ad Dauid et ego custodem capitis mei ponam te cunctis diebus 28:3 Samuhel autem mortuus est planxitque eum 28:3 Now Samuel was dead, and all Israel had lamented

him, and buried him in Ramah, even in his own city. And Saul had put away those that had familiar spirits, and the

omnis Israhel et sepelierunt eum in Rama urbe sua et Saul wizards, out of the land.

abstulit magos et ariolos de terra 28:4 congregatique sunt 28:4 And the Philistines gathered themselves together, and came and pitched in Shunem: and Saul gathered all Israel together, and they pitched in Gilboa.

Philisthim et uenerunt et castrametati sunt in Sunam congregauit autem et Saul uniuersum Israhel et uenit in Gelboe 28:5 et uidit Saul castra Philisthim et timuit et expauit cor 28:5 And when Saul saw the host of the Philistines, he was

afraid, and his heart greatly trembled.

Biblia Sacra Vulgata

28:6 And when Saul enquired of the LORD, the LORD eius nimis 28:6 consuluitque Dominum et non respondit ei

answered him not, neither by dreams, nor by Urim, nor by prophets.

neque per somnia neque per sacerdotes neque per prophe-

28:7 Then said Saul unto his servants, Seek me a woman tas 28:7 dixitque Saul seruis suis quaerite mihi mulierem

that hath a familiar spirit, that I may go to her, and enquire of her. And his servants said to him, Behold, there is a

woman that hath a familiar spirit at Endor. habentem pythonem et uadam ad eam et sciscitabor per il-

lam et dixerunt serui eius ad eum est mulier habens pytho-

28:8 And Saul disguised himself, and put on other raiment, nem in Aendor 28:8 mutauit ergo habitum suum uestitusand he went, and two men with him, and they came to the woman by night: and he said, I pray thee, divine unto me by the familiar spirit, and bring me him up, whom I shall name unto thee.

que est aliis uestimentis abiit ipse et duo uiri cum eo ueneruntque ad mulierem nocte et ait diuina mihi in pythone et

28:9 And the woman said unto him, Behold, thou knowest suscita mihi quem dixero tibi 28:9 et ait mulier ad eum ecce

what Saul hath done, how he hath cut off those that have familiar spirits, and the wizards, out of the land: wherefore

then layest thou a snare for my life, to cause me to die? tu nosti quanta fecerit Saul et quomodo eraserit magos et

ariolos de terra quare ergo insidiaris animae meae ut occidar

28:10 And Saul sware to her by the LORD, saying, As the 28:10 et iurauit ei Saul in Domino dicens uiuit Dominus quia

LORD liveth, there shall no punishment happen to thee for this thing.

non ueniet tibi quicquam mali propter hanc rem 28:11 dixit-

28:11 Then said the woman, Whom shall I bring up unto thee? And he said, Bring me up Samuel.

que ei mulier quem suscitabo tibi qui ait Samuhelem suscita

28:12 And when the woman saw Samuel, she cried with mihi 28:12 cum autem uidisset mulier Samuhelem excla-

a loud voice: and the woman spake to Saul, saying, Why hast thou deceived me? for thou art Saul.

mauit uoce magna et dixit ad Saul quare inposuisti mihi tu

28:13 And the king said unto her, Be not afraid: for what es enim Saul 28:13 dixitque ei rex noli timere quid uidisti

sawest thou? And the woman said unto Saul, I saw gods ascending out of the earth.

et ait mulier ad Saul deos uidi ascendentes de terra 28:14

28:14 And he said unto her, What form is he of? And she said, An old man cometh up; and he is covered with a mantle. And Saul perceived that it was Samuel, and he stooped with his face to the ground, and bowed himself.

dixitque ei qualis est forma eius quae ait uir senex ascendit et ipse amictus est pallio intellexit Saul quod Samuhel esset

28:15 And Samuel said to Saul, Why hast thou disquieted et inclinauit se super faciem suam in terra et adorauit 28:15

me, to bring me up? And Saul answered, I am sore distressed; for the Philistines make war against me, and God is departed from me, and answereth me no more, neither by prophets, nor by dreams: therefore I have called thee, that thou mayest make known unto me what I shall do.

dixit autem Samuhel ad Saul quare inquietasti me ut suscitarer et ait Saul coartor nimis siquidem Philisthim pugnant aduersum me et Deus recessit a me et exaudire me noluit neque in manu prophetarum neque per somnia uocaui ergo

28:16 Then said Samuel, Wherefore then dost thou ask of te ut ostenderes mihi quid faciam 28:16 et ait Samuhel quid

me, seeing the LORD is departed from thee, and is become thine enemy?

interrogas me cum Dominus recesserit a te et transierit ad

28:17 And the LORD hath done to him, as he spake by me: aemulum tuum 28:17 faciet enim Dominus tibi sicut locu-

for the LORD hath rent the kingdom out of thine hand, and given it to thy neighbour, even to David:

tus est in manu mea et scindet regnum de manu tua et da-

28:18 Because thou obeyedst not the voice of the LORD, bit illud proximo tuo Dauid 28:18 quia non oboedisti uoci

nor executedst his fierce wrath upon Amalek, therefore hath the LORD done this thing unto thee this day.

Domini neque fecisti iram furoris eius in Amalech idcirco

28:19 Moreover the LORD will also deliver Israel with quod pateris fecit tibi Dominus hodie 28:19 et dabit Domi-

thee into the hand of the Philistines: and to morrow shalt thou and thy sons be with me: the LORD also shall deliver

the host of Israel into the hand of the Philistines. nus etiam Israhel tecum in manu Philisthim cras autem tu et

Liber I Samuelis

filii tui mecum eritis sed et castra Israhel tradet Dominus in manu Philisthim 28:20 statimque Saul cecidit porrectus in 28:20 Then Saul fell straightway all along on the earth,

and was sore afraid, because of the words of Samuel: and there was no strength in him; for he had eaten no bread all

terram extimuerat enim uerba Samuhel et robur non erat in the day, nor all the night. eo quia non comederat panem tota die illa 28:21 ingressa 28:21 And the woman came unto Saul, and saw that he

was sore troubled, and said unto him, Behold, thine handmaid hath obeyed thy voice, and I have put my life in my hand, and have hearkened unto thy words which thou spakest unto me.

est itaque mulier ad Saul et ait conturbatus enim erat ualde dixitque ad eum ecce oboediuit ancilla tua uoci tuae et posui animam meam in manu mea et audiui sermones tuos quos locutus es ad me 28:22 nunc igitur audi et tu uocem ancillae 28:22 Now therefore, I pray thee, hearken thou also unto

the voice of thine handmaid, and let me set a morsel of bread before thee; and eat, that thou mayest have strength,

tuae ut ponam coram te buccellam panis et comedens con- when thou goest on thy way. ualescas ut possis iter facere 28:23 qui rennuit et ait non 28:23 But he refused, and said, I will not eat. But his

servants, together with the woman, compelled him; and he hearkened unto their voice. So he arose from the earth,

comedam coegerunt autem eum serui sui et mulier et tan- and sat upon the bed. dem audita uoce eorum surrexit de terra et sedit super lectum 28:24 mulier autem illa habebat uitulum pascualem in domo 28:24 And the woman had a fat calf in the house; and she

hasted, and killed it, and took flour, and kneaded it, and did bake unleavened bread thereof:

et festinauit et occidit eum tollensque farinam miscuit eam et coxit azyma 28:25 et posuit ante Saul et ante seruos eius 28:25 And she brought it before Saul, and before his ser-

vants; and they did eat. Then they rose up, and went away that night.

qui cum comedissent surrexerunt et ambulauerunt per totam noctem illam

29:1 congregata sunt ergo Philisthim uniuersa agmina in 29:1 Now the Philistines gathered together all their armies to Aphek: and the Israelites pitched by a fountain which is in Jezreel.

Afec sed et Israhel castrametatus est super fontem qui erat in Iezrahel 29:2 et satrapae quidem Philisthim incedebant 29:2 And the lords of the Philistines passed on by hun-

dreds, and by thousands: but David and his men passed on in the rereward with Achish.

in centuriis et milibus Dauid autem et uiri eius erant in nouissimo agmine cum Achis 29:3 dixeruntque principes Phi- 29:3 Then said the princes of the Philistines, What do

these Hebrews here? And Achish said unto the princes of the Philistines, Is not this David, the servant of Saul the king of Israel, which hath been with me these days, or these years, and I have found no fault in him since he fell unto me unto this day?

listhim quid sibi uolunt Hebraei isti et ait Achis ad principes Philisthim num ignoratis Dauid qui fuit seruus Saul regis Israhel et est apud me multis diebus uel annis et non inueni in eo quicquam ex die qua transfugit ad me usque ad diem hanc 29:4 irati sunt autem aduersus eum principes Philisthim et 29:4 And the princes of the Philistines were wroth with

him; and the princes of the Philistines said unto him, Make this fellow return, that he may go again to his place which thou hast appointed him, and let him not go down with us to battle, lest in the battle he be an adversary to us: for wherewith should he reconcile himself unto his master? should it not be with the heads of these men?

dixerunt ei reuertatur uir et sedeat in loco suo in quo constituisti eum et non descendat nobiscum in proelium ne fiat nobis aduersarius cum proeliari coeperimus quomodo enim aliter placare poterit dominum suum nisi in capitibus nostris 29:5 nonne iste est Dauid cui cantabant in choro dicentes per- 29:5 Is not this David, of whom they sang one to another

in dances, saying, Saul slew his thousands, and David his ten thousands?

cussit Saul in milibus suis et Dauid in decem milibus suis

Biblia Sacra Vulgata

29:6 Then Achish called David, and said unto him, Surely, 29:6 uocauit ergo Achis Dauid et ait ei uiuit Dominus quia as the LORD liveth, thou hast been upright, and thy going out and thy coming in with me in the host is good in my sight: for I have not found evil in thee since the day of thy coming unto me unto this day: nevertheless the lords favour thee not.

rectus es tu et bonus in conspectu meo et exitus tuus et introitus tuus mecum est in castris et non inueni in te quicquam mali ex die qua uenisti ad me usque ad diem hanc sed sat-

29:7 Wherefore now return, and go in peace, that thou dis- rapis non places 29:7 reuertere ergo et uade in pace et non please not the lords of the Philistines.

29:8 And David said unto Achish, But what have I done? offendes oculos satraparum Philisthim 29:8 dixitque Dauid

and what hast thou found in thy servant so long as I have been with thee unto this day, that I may not go fight against

the enemies of my lord the king? ad Achis quid enim feci et quid inuenisti in me seruo tuo

a die qua fui in conspectu tuo usque in diem hanc ut non

29:9 And Achish answered and said to David, I know that ueniam et pugnem contra inimicos domini mei regis 29:9

thou art good in my sight, as an angel of God: notwithstanding the princes of the Philistines have said, He shall

not go up with us to the battle. respondens autem Achis locutus est ad Dauid scio quia bo-

nus es tu in oculis meis sicut angelus Dei sed principes Phi-

29:10 Wherefore now rise up early in the morning with listhim dixerunt non ascendet nobiscum in proelium 29:10

thy master’s servants that are come with thee: and as soon as ye be up early in the morning, and have light, depart.

igitur consurge mane tu et serui domini tui qui uenerunt tecum et cum de nocte surrexeritis et coeperit dilucescere per-

29:11 So David and his men rose up early to depart in the gite 29:11 surrexit itaque de nocte Dauid ipse et uiri eius ut

morning, to return into the land of the Philistines. And the Philistines went up to Jezreel.

proficiscerentur mane et reuerterentur ad terram Philisthim Philisthim autem ascenderunt in Iezrahel

30:1 And it came to pass, when David and his men were 30:1 cumque uenissent Dauid et uiri eius in Siceleg die ter-

come to Ziklag on the third day, that the Amalekites had invaded the south, and Ziklag, and smitten Ziklag, and

burned it with fire; tia Amalechitae impetum fecerant ex parte australi in Sice-30:2 And had taken the women captives, that were therein: leg et percusserant Siceleg et succenderant eam igni 30:2 et

they slew not any, either great or small, but carried them away, and went on their way.

captiuas duxerant mulieres ex ea et a minimo usque ad magnum et non interfecerant quemquam sed secum duxerant et

30:3 So David and his men came to the city, and, behold, it pergebant in itinere suo 30:3 cum ergo uenisset Dauid et was burned with fire; and their wives, and their sons, and their daughters, were taken captives.

uiri eius ad ciuitatem et inuenissent eam succensam igni et

30:4 Then David and the people that were with him lifted uxores suas et filios suos et filias ductas esse captiuas 30:4

up their voice and wept, until they had no more power to weep.

leuauerunt Dauid et populus qui erat cum eo uoces suas et

30:5 And David’s two wives were taken captives, Ahi- planxerunt donec deficerent in eis lacrimae 30:5 siquidem

noam the Jezreelitess, and Abigail the wife of Nabal the Carmelite.

et duae uxores Dauid captiuae ductae fuerant Ahinoem Iez-

30:6 And David was greatly distressed; for the people rahelites et Abigail uxor Nabal Carmeli 30:6 et contrista-

spake of stoning him, because the soul of all the people was grieved, every man for his sons and for his daughters:

but David encouraged himself in the LORD his God. tus est Dauid ualde uolebat enim eum populus lapidare quia

amara erat anima uniuscuiusque uiri super filiis suis et filia-

30:7 And David said to Abiathar the priest, Ahimelech’s bus confortatus est autem Dauid in Domino Deo suo 30:7 et son, I pray thee, bring me hither the ephod. And Abiathar brought thither the ephod to David.

ait ad Abiathar sacerdotem filium Ahimelech adplica ad me

Liber I Samuelis

ephod et adplicuit Abiathar ephod ad Dauid 30:8 et con- 30:8 And David enquired at the LORD, saying, Shall I

pursue after this troop? shall I overtake them? And he answered him, Pursue: for thou shalt surely overtake them,

suluit Dauid Dominum dicens persequar an non latrunculos and without fail recover all. hos et conprehendam eos dixitque ei persequere absque dubio enim conprehendes eos et excuties praedam 30:9 abiit 30:9 So David went, he and the six hundred men that were

with him, and came to the brook Besor, where those that were left behind stayed.

ergo Dauid ipse et sescenti uiri qui erant cum eo et uenerunt usque ad torrentem Besor et lassi quidam substiterunt 30:10 30:10 But David pursued, he and four hundred men: for

two hundred abode behind, which were so faint that they could not go over the brook Besor.

persecutus est autem Dauid ipse et quadringenti uiri substiterant enim ducenti qui lassi transire non poterant torrentem Besor 30:11 et inuenerunt uirum aegyptium in agro et ad- 30:11 And they found an Egyptian in the field, and brought

him to David, and gave him bread, and he did eat; and they made him drink water;

duxerunt eum ad Dauid dederuntque ei panem ut comederet et ut biberet aquam 30:12 sed et fragmen massae caricarum 30:12 And they gave him a piece of a cake of figs, and two

clusters of raisins: and when he had eaten, his spirit came again to him: for he had eaten no bread, nor drunk any

et duas ligaturas uuae passae quae cum comedisset reuer- water, three days and three nights. sus est spiritus eius et refocilatus est non enim comederat panem neque biberat aquam tribus diebus et tribus noctibus 30:13 dixit itaque ei Dauid cuius es tu uel unde quo pergis qui 30:13 And David said unto him, To whom belongest thou?

and whence art thou? And he said, I am a young man of Egypt, servant to an Amalekite; and my master left me,

ait ei puer aegyptius ego sum seruus uiri amalechitae dere- because three days agone I fell sick. liquit autem me dominus meus quia aegrotare coepi nudius tertius 30:14 siquidem nos erupimus ad australem partem 30:14 We made an invasion upon the south of the

Cherethites, and upon the coast which belongeth to Judah, and upon the south of Caleb; and we burned Ziklag with

Cerethi et contra Iudam et ad meridiem Chaleb et Siceleg fire. succendimus igni 30:15 dixitque ei Dauid potes me ducere 30:15 And David said to him, Canst thou bring me down

to this company? And he said, Swear unto me by God, that thou wilt neither kill me, nor deliver me into the hands of

ad istum cuneum qui ait iura mihi per Deum quod non occi- my master, and I will bring thee down to this company. das me et non tradas me in manu domini mei et ducam te ad cuneum istum 30:16 qui cum duxisset eum ecce illi discum- 30:16 And when he had brought him down, behold, they

were spread abroad upon all the earth, eating and drinking, and dancing, because of all the great spoil that they had taken out of the land of the Philistines, and out of the land of Judah.

bebant super faciem uniuersae terrae comedentes et bibentes et quasi festum celebrantes diem pro cuncta praeda et spoliis quae ceperant de terra Philisthim et de terra Iuda 30:17 30:17 And David smote them from the twilight even unto

the evening of the next day: and there escaped not a man of them, save four hundred young men, which rode upon

et percussit eos Dauid a uespere usque ad uesperam alterius camels, and fled. diei et non euasit ex eis quisquam nisi quadringenti uiri adulescentes qui ascenderant camelos et fugerant 30:18 eruit 30:18 And David recovered all that the Amalekites had

carried away: and David rescued his two wives.

ergo Dauid omnia quae tulerant Amalechitae et duas uxores suas eruit 30:19 nec defuit quicquam a paruo usque ad ma- 30:19 And there was nothing lacking to them, neither

small nor great, neither sons nor daughters, neither spoil, nor any thing that they had taken to them: David recovered

gnum tam de filiis quam de filiabus et de spoliis et quaecum- all. que rapuerant omnia reduxit Dauid 30:20 et tulit uniuersos 30:20 And David took all the flocks and the herds, which

they drave before those other cattle, and said, This is David’s spoil.

Biblia Sacra Vulgata

greges et armenta et minauit ante faciem suam dixeruntque

30:21 And David came to the two hundred men, which haec est praeda Dauid 30:21 uenit autem Dauid ad ducentos

were so faint that they could not follow David, whom they had made also to abide at the brook Besor: and they went forth to meet David, and to meet the people that were with him: and when David came near to the people, he saluted them.

uiros qui lassi substiterant nec sequi potuerant Dauid et residere eos iusserat in torrente Besor qui egressi sunt obuiam Dauid et populo qui erat cum eo accedens autem Dauid ad

30:22 Then answered all the wicked men and men of Be- populum salutauit eos pacifice 30:22 respondensque omnis

lial, of those that went with David, and said, Because they went not with us, we will not give them ought of the spoil that we have recovered, save to every man his wife and his children, that they may lead them away, and depart.

uir pessimus et iniquus de uiris qui ierant cum Dauid dixit quia non uenerunt nobiscum non dabimus eis quicquam de praeda quam eruimus sed sufficiat unicuique uxor sua et filii

30:23 Then said David, Ye shall not do so, my brethren, quos cum acceperint recedant 30:23 dixit autem Dauid non

with that which the LORD hath given us, who hath preserved us, and delivered the company that came against us

into our hand. sic facietis fratres mei de his quae tradidit Dominus nobis

et custodiuit nos et dedit latrunculos qui eruperant aduersum

30:24 For who will hearken unto you in this matter? but nos in manu nostra 30:24 nec audiet uos quisquam super as his part is that goeth down to the battle, so shall his part be that tarrieth by the stuff: they shall part alike.

sermone hoc aequa enim pars erit descendentis ad proelium

30:25 And it was so from that day forward, that he made et remanentis ad sarcinas et similiter diuident 30:25 et fac-

it a statute and an ordinance for Israel unto this day.

tum est hoc ex die illa et deinceps constitutum et praefinitum

30:26 And when David came to Ziklag, he sent of the spoil et quasi lex in Israhel usque ad diem hanc 30:26 uenit ergo

unto the elders of Judah, even to his friends, saying, Behold a present for you of the spoil of the enemies of the

LORD; Dauid in Siceleg et misit dona de praeda senioribus Iuda pro-

ximis suis dicens accipite benedictionem de praeda hostium

30:27 To them which were in Bethel, and to them which Domini 30:27 his qui erant in Bethel et qui in Ramoth ad

were in south Ramoth, and to them which were in Jattir,

30:28 And to them which were in Aroer, and to them meridiem et qui in Iether 30:28 et qui in Aroer et qui in Se-

which were in Siphmoth, and to them which were in Eshtemoa,

famoth et qui in Esthama 30:29 et qui in Rachal et qui in ur-

30:29 And to them which were in Rachal, and to them which were in the cities of the Jerahmeelites, and to them which were in the cities of the Kenites,

bibus Ierameli et qui in urbibus Ceni 30:30 et qui in Arama

30:30 And to them which were in Hormah, and to them which were in Chorashan, and to them which were in

Athach, et qui in lacu Asan et qui in Athac 30:31 et qui in Hebron

30:31 And to them which were in Hebron, and to all the places where David himself and his men were wont to haunt.

et reliquis qui erant in his locis in quibus commoratus fuerat Dauid ipse et uiri eius

31:1 Now the Philistines fought against Israel: and the 31:1 Philisthim autem pugnabant aduersum Israhel et fu-

men of Israel fled from before the Philistines, and fell down slain in mount Gilboa.

gerunt uiri Israhel ante faciem Philisthim et ceciderunt inter-

31:2 And the Philistines followed hard upon Saul and fecti in monte Gelboe 31:2 inrueruntque Philisthim in Saul

upon his sons; and the Philistines slew Jonathan, and Abinadab, and Melchishua, Saul’s sons.

et filios eius et percusserunt Ionathan et Abinadab et Mel-

31:3 And the battle went sore against Saul, and the archers chisue filios Saul 31:3 totumque pondus proelii uersum est hit him; and he was sore wounded of the archers.

in Saul et consecuti sunt eum uiri sagittarii et uulneratus est

31:4 Then said Saul unto his armourbearer, Draw thy uehementer a sagittariis 31:4 dixitque Saul ad armigerum

sword, and thrust me through therewith; lest these uncircumcised come and thrust me through, and abuse me. But his armourbearer would not; for he was sore afraid. Therefore Saul took a sword, and fell upon it.

Liber I Samuelis

suum euagina gladium tuum et percute me ne forte ueniant incircumcisi isti et interficiant me inludentes mihi et noluit armiger eius fuerat enim nimio timore perterritus arripuit ita-

que Saul gladium et inruit super eum 31:5 quod cum ui- 31:5 And when his armourbearer saw that Saul was dead,

he fell likewise upon his sword, and died with him.

disset armiger eius uidelicet quod mortuus esset Saul inruit etiam ipse super gladium suum et mortuus est cum eo 31:6 31:6 So Saul died, and his three sons, and his ar-

mourbearer, and all his men, that same day together.

mortuus est ergo Saul et tres filii eius et armiger illius et uniuersi uiri eius in die illa pariter 31:7 uidentes autem uiri 31:7 And when the men of Israel that were on the other

side of the valley, and they that were on the other side Jordan, saw that the men of Israel fled, and that Saul and his sons were dead, they forsook the cities, and fled; and the Philistines came and dwelt in them.

Israhel qui erant trans uallem et trans Iordanem quod fugissent uiri israhelitae et quod mortuus esset Saul et filii eius reliquerunt ciuitates suas et fugerunt ueneruntque Philisthim et habitauerunt ibi 31:8 facta autem die altera uenerunt Phi- 31:8 And it came to pass on the morrow, when the

Philistines came to strip the slain, that they found Saul and his three sons fallen in mount Gilboa.

listhim ut spoliarent interfectos et inuenerunt Saul et tres filios eius iacentes in monte Gelboe 31:9 et praeciderunt ca- 31:9 And they cut off his head, and stripped off his armour,

and sent into the land of the Philistines round about, to publish it in the house of their idols, and among the people.

put Saul et expoliauerunt eum armis et miserunt in terram Philisthinorum per circuitum ut adnuntiaretur in templo idolorum et in populis 31:10 et posuerunt arma eius in templo 31:10 And they put his armour in the house of Ashtaroth:

and they fastened his body to the wall of Bethshan.

Astharoth corpus uero eius suspenderunt in muro Bethsan 31:11 quod cum audissent habitatores Iabesgalaad quaecum- 31:11 And when the inhabitants of Jabeshgilead heard of

that which the Philistines had done to Saul;

que fecerant Philisthim Saul 31:12 surrexerunt omnes uiri 31:12 All the valiant men arose, and went all night, and

took the body of Saul and the bodies of his sons from the wall of Bethshan, and came to Jabesh, and burnt them

fortissimi et ambulauerunt tota nocte et tulerunt cadauer Saul there. et cadauera filiorum eius de muro Bethsan ueneruntque Iabes et conbuserunt ea ibi 31:13 et tulerunt ossa eorum et 31:13 And they took their bones, and buried them under a

tree at Jabesh, and fasted seven days.

sepelierunt in nemore Iabes et ieiunauerunt septem diebus

Biblia Sacra Vulgata

DXII

Liber II Samuelis

1:1 factum est autem postquam mortuus est Saul ut Dauid 1:1 Now it came to pass after the death of Saul, when

David was returned from the slaughter of the Amalekites, and David had abode two days in Ziklag;

reuerteretur a caede Amalech et maneret in Siceleg dies duos 1:2 in die autem tertia apparuit homo ueniens de castris Saul 1:2 It came even to pass on the third day, that, behold, a

man came out of the camp from Saul with his clothes rent, and earth upon his head: and so it was, when he came to

ueste conscissa et puluere aspersus caput et ut uenit ad Da- David, that he fell to the earth, and did obeisance. uid cecidit super faciem suam et adorauit 1:3 dixitque ad 1:3 And David said unto him, From whence comest thou?

And he said unto him, Out of the camp of Israel am I escaped.

eum Dauid unde uenis qui ait ad eum de castris Israhel fugi 1:4 et dixit ad eum Dauid quod est uerbum quod factum est 1:4 And David said unto him, How went the matter? I pray

thee, tell me. And he answered, That the people are fled from the battle, and many of the people also are fallen and

indica mihi qui ait fugit populus e proelio et multi corruentes dead; and Saul and Jonathan his son are dead also. e populo mortui sunt sed et Saul et Ionathan filius eius interierunt 1:5 dixitque Dauid ad adulescentem qui nuntiabat 1:5 And David said unto the young man that told him,

How knowest thou that Saul and Jonathan his son be dead?

unde scis quia mortuus est Saul et Ionathan filius eius 1:6 1:6 And the young man that told him said, As I happened

by chance upon mount Gilboa, behold, Saul leaned upon his spear; and, lo, the chariots and horsemen followed hard

ait adulescens qui narrabat ei casu ueni in montem Gelboe et after him. Saul incumbebat super hastam suam porro currus et equites adpropinquabant ei 1:7 et conuersus post tergum suum ui- 1:7 And when he looked behind him, he saw me, and

called unto me. And I answered, Here am I.

densque me uocauit cui cum respondissem adsum 1:8 dixit 1:8 And he said unto me, Who art thou? And I answered

him, I am an Amalekite.

mihi quisnam es tu et aio ad eum Amalechites sum 1:9 et 1:9 He said unto me again, Stand, I pray thee, upon me,

and slay me: for anguish is come upon me, because my life is yet whole in me.

locutus est mihi sta super me et interfice me quoniam tenent me angustiae et adhuc tota anima in me est 1:10 stansque 1:10 So I stood upon him, and slew him, because I was

sure that he could not live after that he was fallen: and I took the crown that was upon his head, and the bracelet that was on his arm, and have brought them hither unto my lord.

super eum occidi illum sciebam enim quod uiuere non poterat post ruinam et tuli diadema quod erat in capite eius et armillam de brachio illius et adtuli ad te dominum meum huc 1:11 adprehendens autem Dauid uestimenta sua scidit 1:11 Then David took hold on his clothes, and rent them;

and likewise all the men that were with him:

omnesque uiri qui erant cum eo 1:12 et planxerunt et fle- 1:12 And they mourned, and wept, and fasted until even,

for Saul, and for Jonathan his son, and for the people of the LORD, and for the house of Israel; because they were

uerunt et ieiunauerunt usque ad uesperam super Saul et su- fallen by the sword. per Ionathan filium eius et super populum Domini et super domum Israhel quod corruissent gladio 1:13 dixitque Da- 1:13 And David said unto the young man that told him,

Whence art thou? And he answered, I am the son of a stranger, an Amalekite.

uid ad iuuenem qui nuntiauerat ei unde es qui respondit filius hominis aduenae amalechitae ego sum 1:14 et ait ad 1:14 And David said unto him, How wast thou not

afraid to stretch forth thine hand to destroy the LORD’s anointed?

eum Dauid quare non timuisti mittere manum tuam ut occideres christum Domini 1:15 uocansque Dauid unum de 1:15 And David called one of the young men, and said, Go

near, and fall upon him. And he smote him that he died.

pueris ait accedens inrue in eum qui percussit illum et mortuus est 1:16 et ait ad eum Dauid sanguis tuus super caput 1:16 And David said unto him, Thy blood be upon thy

head; for thy mouth hath testified against thee, saying, I have slain the LORD’s anointed.

tuum os enim tuum locutum est aduersum te dicens ego in-

Biblia Sacra Vulgata

1:17 And David lamented with this lamentation over Saul terfeci christum Domini 1:17 planxit autem Dauid planc-

and over Jonathan his son:

tum huiuscemodi super Saul et super Ionathan filium eius

1:18 (Also he bade them teach the children of Judah the 1:18 et praecepit ut docerent filios Iuda arcum sicut scrip-

use of the bow: behold, it is written in the book of Jasher.)

1:19 The beauty of Israel is slain upon thy high places: tum est in libro Iustorum 1:19 incliti Israhel super montes

how are the mighty fallen!

1:20 Tell it not in Gath, publish it not in the streets of tuos interfecti sunt quomodo ceciderunt fortes 1:20 nolite

Askelon; lest the daughters of the Philistines rejoice, lest the daughters of the uncircumcised triumph.

adnuntiare in Geth neque adnuntietis in conpetis Ascalonis ne forte laetentur filiae Philisthim ne exultent filiae incir-

1:21 Ye mountains of Gilboa, let there be no dew, neither cumcisorum 1:21 montes Gelboe nec ros nec pluuiae uelet there be rain, upon you, nor fields of offerings: for there the shield of the mighty is vilely cast away, the shield of

Saul, as though he had not been anointed with oil. niant super uos neque sint agri primitiarum quia ibi abiec-

tus est clypeus fortium clypeus Saul quasi non esset unctus

1:22 From the blood of the slain, from the fat of the oleo 1:22 a sanguine interfectorum ab adipe fortium sagitta

mighty, the bow of Jonathan turned not back, and the sword of Saul returned not empty.

Ionathan numquam rediit retrorsum et gladius Saul non est

1:23 Saul and Jonathan were lovely and pleasant in their reuersus inanis 1:23 Saul et Ionathan amabiles et decori in

lives, and in their death they were not divided: they were swifter than eagles, they were stronger than lions.

uita sua in morte quoque non sunt diuisi aquilis uelociores

1:24 Ye daughters of Israel, weep over Saul, who clothed leonibus fortiores 1:24 filiae Israhel super Saul flete qui ue-

you in scarlet, with other delights, who put on ornaments of gold upon your apparel.

stiebat uos coccino in deliciis qui praebebat ornamenta au-

1:25 How are the mighty fallen in the midst of the battle! rea cultui uestro 1:25 quomodo ceciderunt fortes in proelio

O Jonathan, thou wast slain in thine high places.

1:26 I am distressed for thee, my brother Jonathan: very Ionathan in excelsis tuis occisus est 1:26 doleo super te fra-

pleasant hast thou been unto me: thy love to me was wonderful, passing the love of women.

ter mi Ionathan decore nimis et amabilis super amorem mu-

1:27 How are the mighty fallen, and the weapons of war lierum 1:27 quomodo ceciderunt robusti et perierunt arma

perished!

bellica

2:1 And it came to pass after this, that David enquired of 2:1 igitur post haec consuluit Dauid Dominum dicens num the LORD, saying, Shall I go up into any of the cities of Judah? And the LORD said unto him, Go up. And David

said, Whither shall I go up? And he said, Unto Hebron. ascendam in unam de ciuitatibus Iuda et ait Dominus ad eum

ascende dixitque Dauid quo ascendam et respondit ei in He-

2:2 So David went up thither, and his two wives also, bron 2:2 ascendit ergo Dauid et duae uxores eius Ahinoem

Ahinoam the Jezreelitess, and Abigail Nabal’s wife the Carmelite.

Iezrahelites et Abigail uxor Nabal Carmeli 2:3 sed et ui-

2:3 And his men that were with him did David bring up, every man with his household: and they dwelt in the cities of Hebron.

ros qui erant cum eo duxit Dauid singulos cum domo sua

2:4 And the men of Judah came, and there they anointed et manserunt in oppidis Hebron 2:4 ueneruntque uiri Iuda

David king over the house of Judah. And they told David, saying, That the men of Jabeshgilead were they that buried

Saul. et unxerunt ibi Dauid ut regnaret super domum Iuda et nun-2:5 And David sent messengers unto the men of tiatum est Dauid quod uiri Iabesgalaad sepelissent Saul 2:5

Jabeshgilead, and said unto them, Blessed be ye of the LORD, that ye have shewed this kindness unto your lord,

even unto Saul, and have buried him. misit ergo Dauid nuntios ad uiros Iabesgalaad dixitque ad

eos benedicti uos Domino qui fecistis misericordiam hanc

2:6 And now the LORD shew kindness and truth unto you: cum domino uestro Saul et sepelistis eum 2:6 et nunc re-

and I also will requite you this kindness, because ye have done this thing.

Liber II Samuelis

tribuet quidem uobis Dominus misericordiam et ueritatem sed et ego reddam gratiam eo quod feceritis uerbum istud 2:7 confortentur manus uestrae et estote filii fortitudinis li- 2:7 Therefore now let your hands be strengthened, and be

ye valiant: for your master Saul is dead, and also the house of Judah have anointed me king over them.

cet enim mortuus sit dominus uester Saul tamen me unxit

domus Iuda regem sibi 2:8 Abner autem filius Ner prin- 2:8 But Abner the son of Ner, captain of Saul’s host, took Ishbosheth the son of Saul, and brought him over to Mahanaim;

ceps exercitus Saul tulit Hisboseth filium Saul et circumduxit eum per Castra 2:9 regemque constituit super Galaad 2:9 And made him king over Gilead, and over the Ashu-

rites, and over Jezreel, and over Ephraim, and over Benjamin, and over all Israel.

et super Gesuri et super Iezrahel et super Ephraim et super Beniamin et super Israhel uniuersum 2:10 quadraginta an- 2:10 Ishbosheth Saul’s son was forty years old when he

began to reign over Israel, and reigned two years. But the house of Judah followed David.

norum erat Hisboseth filius Saul cum regnare coepisset super Israhel et duobus annis regnauit sola autem domus Iuda sequebatur Dauid 2:11 et fuit numerus dierum quos com- 2:11 And the time that David was king in Hebron over the

house of Judah was seven years and six months.

moratus est Dauid imperans in Hebron super domum Iuda septem annorum et sex mensuum 2:12 egressusque Abner 2:12 And Abner the son of Ner, and the servants of Ish-

bosheth the son of Saul, went out from Mahanaim to Gibeon.

filius Ner et pueri Hisboseth filii Saul de Castris in Gabaon 2:13 porro Ioab filius Saruiae et pueri Dauid egressi sunt et 2:13 And Joab the son of Zeruiah, and the servants of

David, went out, and met together by the pool of Gibeon: and they sat down, the one on the one side of the pool, and

occurrerunt eis iuxta piscinam Gabaon et cum in unum con- the other on the other side of the pool. uenissent e regione sederunt hii ex una parte piscinae et illi ex altera 2:14 dixitque Abner ad Ioab surgant pueri et lu- 2:14 And Abner said to Joab, Let the young men now

arise, and play before us. And Joab said, Let them arise.

dant coram nobis et respondit Ioab surgant 2:15 surrexe- 2:15 Then there arose and went over by number twelve of

Benjamin, which pertained to Ishbosheth the son of Saul, and twelve of the servants of David.

runt ergo et transierunt numero duodecim de Beniamin ex parte Hisboseth filii Saul et duodecim de pueris Dauid 2:16 2:16 And they caught every one his fellow by the head, and

thrust his sword in his fellow’s side; so they fell down together: wherefore that place was called Helkathhazzurim,

adprehensoque unusquisque capite conparis sui defixit gla- which is in Gibeon. dium in latus contrarii et ceciderunt simul uocatumque est nomen loci illius ager Robustorum in Gabaon 2:17 et or- 2:17 And there was a very sore battle that day; and Abner

was beaten, and the men of Israel, before the servants of David.

tum est bellum durum satis in die illa fugatusque est Abner et uiri Israhel a pueris Dauid 2:18 erant autem ibi tres filii 2:18 And there were three sons of Zeruiah there, Joab, and

Abishai, and Asahel: and Asahel was as light of foot as a wild roe.

Saruiae Ioab et Abisai et Asahel porro Asahel cursor uelocissimus fuit quasi unus ex capreis quae morantur in siluis 2:19 persequebatur autem Asahel Abner et non declinauit ad 2:19 And Asahel pursued after Abner; and in going he

turned not to the right hand nor to the left from following Abner.

dexteram siue ad sinistram omittens persequi Abner 2:20

2:20 Then Abner looked behind him, and said, Art thou Asahel? And he answered, I am.

respexit itaque Abner post tergum suum et ait tune es Asahel qui respondit ego sum 2:21 dixitque ei Abner uade ad dext- 2:21 And Abner said to him, Turn thee aside to thy right

hand or to thy left, and lay thee hold on one of the young men, and take thee his armour. But Asahel would not turn aside from following of him.

Biblia Sacra Vulgata

ram siue ad sinistram et adprehende unum de adulescentibus et tolle tibi spolia eius noluit autem Asahel omittere quin

2:22 And Abner said again to Asahel, Turn thee aside urgueret eum 2:22 rursumque locutus est Abner ad Asahel

from following me: wherefore should I smite thee to the ground? how then should I hold up my face to Joab thy

brother? recede noli me sequi ne conpellar confodere te in terra et le-

2:23 Howbeit he refused to turn aside: wherefore Abner uare non potero faciem meam ad Ioab fratrem tuum 2:23 with the hinder end of the spear smote him under the fifth rib, that the spear came out behind him; and he fell down there, and died in the same place: and it came to pass, that as many as came to the place where Asahel fell down and died stood still.

qui audire contempsit et noluit declinare percussit ergo eum Abner auersa hasta in inguine et transfodit et mortuus est in eodem loco omnesque qui transiebant per locum in quo ceci-

2:24 Joab also and Abishai pursued after Abner: and derat Asahel et mortuus erat subsistebant 2:24 persequen-

the sun went down when they were come to the hill of Ammah, that lieth before Giah by the way of the wilder-

ness of Gibeon. tibus autem Ioab et Abisai fugientem Abner sol occubuit et

uenerunt usque ad collem Aquaeductus qui est ex aduerso

2:25 And the children of Benjamin gathered themselves uallis et itineris deserti in Gabaon 2:25 congregatique sunt

together after Abner, and became one troop, and stood on the top of an hill.

filii Beniamin ad Abner et conglobati in unum cuneum ste-

2:26 Then Abner called to Joab, and said, Shall the sword terunt in summitate tumuli unius 2:26 et exclamauit Abner

devour for ever? knowest thou not that it will be bitterness in the latter end? how long shall it be then, ere thou bid

the people return from following their brethren? ad Ioab et ait num usque ad internicionem tuus mucro desae-

uiet an ignoras quod periculosa sit desperatio usquequo non

2:27 And Joab said, As God liveth, unless thou hadst spo- dicis populo ut omittat persequi fratres suos 2:27 et ait Ioab

ken, surely then in the morning the people had gone up every one from following his brother.

uiuit Dominus si locutus fuisses mane recessisset populus

2:28 So Joab blew a trumpet, and all the people stood still, persequens fratrem suum 2:28 insonuit ergo Ioab bucina et and pursued after Israel no more, neither fought they any more.

stetit omnis exercitus nec persecuti sunt ultra Israhel neque

2:29 And Abner and his men walked all that night through iniere certamen 2:29 Abner autem et uiri eius abierunt per

the plain, and passed over Jordan, and went through all Bithron, and they came to Mahanaim.

campestria tota nocte illa et transierunt Iordanem et lustrata

2:30 And Joab returned from following Abner: and when omni Bethoron uenerunt ad Castra 2:30 porro Ioab reuer-

he had gathered all the people together, there lacked of David’s servants nineteen men and Asahel.

sus omisso Abner congregauit omnem populum et defuerunt

2:31 But the servants of David had smitten of Benjamin, de pueris Dauid decem et nouem uiri excepto Asahele 2:31

and of Abner’s men, so that three hundred and threescore men died.

serui autem Dauid percusserunt de Beniamin et de uiris qui

2:32 And they took up Asahel, and buried him in the sepul- erant cum Abner trecentos sexaginta qui et mortui sunt 2:32

chre of his father, which was in Bethlehem. And Joab and his men went all night, and they came to Hebron at break

of day. tuleruntque Asahel et sepelierunt eum in sepulchro patris sui in Bethleem et ambulauerunt tota nocte Ioab et uiri qui erant cum eo et in ipso crepusculo peruenerunt in Hebron

3:1 Now there was long war between the house of Saul 3:1 facta est ergo longa concertatio inter domum Saul et

and the house of David: but David waxed stronger and stronger, and the house of Saul waxed weaker and weaker.

inter domum Dauid Dauid proficiens et semper se ipso robu-

3:2 And unto David were sons born in Hebron: and his stior domus autem Saul decrescens cotidie 3:2 nati quoque

firstborn was Amnon, of Ahinoam the Jezreelitess;

Liber II Samuelis

sunt filii Dauid in Hebron fuitque primogenitus eius Amnon de Ahinoem Iezrahelitide 3:3 et post eum Chelaab de Abi- 3:3 And his second, Chileab, of Abigail the wife of Nabal

the Carmelite; and the third, Absalom the son of Maacah the daughter of Talmai king of Geshur;

gail uxore Nabal Carmeli porro tertius Absalom filius Maacha filiae Tholomai regis Gessur 3:4 quartus autem Ad- 3:4 And the fourth, Adonijah the son of Haggith; and the

fifth, Shephatiah the son of Abital;

onias filius Aggith et quintus Safathia filius Abital 3:5 sex- 3:5 And the sixth, Ithream, by Eglah David’s wife. These

were born to David in Hebron.

tus quoque Iethraam de Agla uxore Dauid hii nati sunt Da-

uid in Hebron 3:6 cum ergo esset proelium inter domum 3:6 And it came to pass, while there was war between the house of Saul and the house of David, that Abner made himself strong for the house of Saul.

Saul et domum Dauid Abner filius Ner regebat domum Saul 3:7 fuerat autem Sauli concubina nomine Respha filia Ahia 3:7 And Saul had a concubine, whose name was Riz-

pah, the daughter of Aiah: and Ishbosheth said to Abner, Wherefore hast thou gone in unto my father’s concubine?

dixitque Hisboseth ad Abner 3:8 quare ingressus es ad con-

3:8 Then was Abner very wroth for the words of Ishbosheth, and said, Am I a dog’s head, which against Judah do shew kindness this day unto the house of Saul thy father, to his brethren, and to his friends, and have not delivered thee into the hand of David, that thou chargest me to day with a fault concerning this woman?

cubinam patris mei qui iratus nimis propter uerba Hisboseth ait numquid caput canis ego sum aduersum Iuda hodie qui fecerim misericordiam super domum Saul patris tui et super fratres et proximos eius et non tradidi te in manu Dauid et tu requisisti in me quod argueres pro muliere hodie 3:9 3:9 So do God to Abner, and more also, except, as the

LORD hath sworn to David, even so I do to him;

haec faciat Deus Abner et haec addat ei nisi quomodo iurauit Dominus Dauid sic faciam cum eo 3:10 ut transferatur re- 3:10 To translate the kingdom from the house of Saul, and

to set up the throne of David over Israel and over Judah, from Dan even to Beersheba.

gnum de domo Saul et eleuetur thronus Dauid super Israhel et super Iudam a Dan usque Bersabee 3:11 et non potuit re- 3:11 And he could not answer Abner a word again, be-

cause he feared him.

spondere ei quicquam quia metuebat illum 3:12 misit ergo 3:12 And Abner sent messengers to David on his behalf,

saying, Whose is the land? saying also, Make thy league with me, and, behold, my hand shall be with thee, to bring

Abner nuntios ad Dauid pro se dicentes cuius est terra et lo- about all Israel unto thee. querentur fac mecum amicitias et erit manus mea tecum et reducam ad te uniuersum Israhel 3:13 qui ait optime ego fa- 3:13 And he said, Well; I will make a league with thee: but

one thing I require of thee, that is, Thou shalt not see my face, except thou first bring Michal Saul’s daughter, when

ciam tecum amicitias sed unam rem peto a te dicens non ui- thou comest to see my face. debis faciem meam antequam adduxeris Michol filiam Saul et sic uenies et uidebis me 3:14 misit autem Dauid nuntios 3:14 And David sent messengers to Ishbosheth Saul’s son,

saying, Deliver me my wife Michal, which I espoused to me for an hundred foreskins of the Philistines.

ad Hisboseth filium Saul dicens redde uxorem meam Michol quam despondi mihi centum praeputiis Philisthim 3:15 mi- 3:15 And Ishbosheth sent, and took her from her husband,

even from Phaltiel the son of Laish.

sit ergo Hisboseth et tulit eam a uiro suo Faltihel filio Lais 3:16 sequebaturque eam uir suus plorans usque Baurim et 3:16 And her husband went with her along weeping be-

hind her to Bahurim. Then said Abner unto him, Go, return. And he returned.

dixit ad eum Abner uade reuertere qui reuersus est 3:17 ser-

3:17 And Abner had communication with the elders of Israel, saying, Ye sought for David in times past to be king over you:

monem quoque intulit Abner ad seniores Israhel dicens tam heri quam nudius tertius quaerebatis Dauid ut regnaret su-

Biblia Sacra Vulgata

3:18 Now then do it: for the LORD hath spoken of David, per uos 3:18 nunc ergo facite quoniam Dominus locutus est

saying, By the hand of my servant David I will save my people Israel out of the hand of the Philistines, and out of

the hand of all their enemies. ad Dauid dicens in manu serui mei Dauid saluabo populum

meum Israhel de manu Philisthim et omnium inimicorum

3:19 And Abner also spake in the ears of Benjamin: and eius 3:19 locutus est autem Abner etiam ad Beniamin et

Abner went also to speak in the ears of David in Hebron all that seemed good to Israel, and that seemed good to the

whole house of Benjamin. abiit ut loqueretur ad Dauid in Hebron omnia quae placue-3:20 So Abner came to David to Hebron, and twenty men rant Israhel et uniuerso Beniamin 3:20 uenitque ad Dauid

with him. And David made Abner and the men that were with him a feast.

in Hebron cum uiginti uiris et fecit Dauid Abner et uiris eius

3:21 And Abner said unto David, I will arise and go, and qui uenerant cum eo conuiuium 3:21 et dixit Abner ad Da-

will gather all Israel unto my lord the king, that they may make a league with thee, and that thou mayest reign over all that thine heart desireth. And David sent Abner away; and he went in peace.

uid surgam ut congregem ad te dominum meum regem omnem Israhel et ineam tecum foedus et imperes omnibus sicut desiderat anima tua cum ergo deduxisset Dauid Abner et

3:22 And, behold, the servants of David and Joab came ille isset in pace 3:22 statim pueri Dauid et Ioab uenerunt

from pursuing a troop, and brought in a great spoil with them: but Abner was not with David in Hebron; for he

had sent him away, and he was gone in peace. caesis latronibus cum praeda magna nimis Abner autem non

erat cum Dauid in Hebron quia iam dimiserat eum et profec-

3:23 When Joab and all the host that was with him were tus fuerat in pace 3:23 et Ioab et omnis exercitus qui erat

come, they told Joab, saying, Abner the son of Ner came to the king, and he hath sent him away, and he is gone in

peace. cum eo postea uenerant nuntiatum est itaque Ioab a narranti-

bus uenit Abner filius Ner ad regem et dimisit eum et abiit in

3:24 Then Joab came to the king, and said, What hast thou pace 3:24 et ingressus est Ioab ad regem et ait quid fecisti

done? behold, Abner came unto thee; why is it that thou hast sent him away, and he is quite gone?

ecce uenit Abner ad te quare dimisisti eum et abiit et reces-

3:25 Thou knowest Abner the son of Ner, that he came to sit 3:25 ignoras Abner filium Ner quoniam ad hoc uenit ut

deceive thee, and to know thy going out and thy coming in, and to know all that thou doest.

deciperet te et sciret exitum tuum et introitum tuum et nos-

3:26 And when Joab was come out from David, he sent set omnia quae agis 3:26 egressus itaque Ioab a Dauid misit

messengers after Abner, which brought him again from the well of Sirah: but David knew it not.

nuntios post Abner et reduxit eum a cisterna Sira ignorante

3:27 And when Abner was returned to Hebron, Joab took Dauid 3:27 cumque rediiset Abner in Hebron seorsum ab-

him aside in the gate to speak with him quietly, and smote him there under the fifth rib, that he died, for the blood of

Asahel his brother. duxit eum Ioab ad medium portae ut loqueretur ei in dolo

et percussit illum ibi in inguine et mortuus est in ultionem

3:28 And afterward when David heard it, he said, I and my sanguinis Asahel fratris eius 3:28 quod cum audisset Dauid

kingdom are guiltless before the LORD for ever from the blood of Abner the son of Ner:

rem iam gestam ait mundus ego sum et regnum meum apud Dominum usque in sempiternum a sanguine Abner filii Ner

3:29 Let it rest on the head of Joab, and on all his father’s 3:29 et ueniat super caput Ioab et super omnem domum pathouse; and let there not fail from the house of Joab one that hath an issue, or that is a leper, or that leaneth on a staff, or that falleth on the sword, or that lacketh bread. ris eius nec deficiat de domo Ioab fluxum seminis sustinens

et leprosus tenens fusum et cadens gladio et indigens pane

3:30 So Joab, and Abishai his brother slew Abner, because 3:30 igitur Ioab et Abisai frater eius interfecerunt Abner eo

he had slain their brother Asahel at Gibeon in the battle.

Liber II Samuelis

quod occidisset Asahel fratrem eorum in Gabaon in proelio 3:31 dixit autem Dauid ad Ioab et ad omnem populum qui 3:31 And David said to Joab, and to all the people that

were with him, Rend your clothes, and gird you with sackcloth, and mourn before Abner. And king David himself

erat cum eo scindite uestimenta uestra et accingimini sac- followed the bier. cis et plangite ante exequias Abner porro rex Dauid sequebatur feretrum 3:32 cumque sepelissent Abner in Hebron 3:32 And they buried Abner in Hebron: and the king lifted

up his voice, and wept at the grave of Abner; and all the people wept.

leuauit rex uocem suam et fleuit super tumulum Abner fleuit autem et omnis populus 3:33 plangensque rex Abner ait 3:33 And the king lamented over Abner, and said, Died

Abner as a fool dieth?

nequaquam ut mori solent ignaui mortuus est Abner 3:34 3:34 Thy hands were not bound, nor thy feet put into fet-

ters: as a man falleth before wicked men, so fellest thou. And all the people wept again over him.

manus tuae non sunt ligatae et pedes tui non sunt conpedibus adgrauati sed sicut solent cadere coram filiis iniquitatis corruisti congeminansque omnis populus fleuit super eum 3:35 cumque uenisset uniuersa multitudo cibum capere cum 3:35 And when all the people came to cause David to eat

meat while it was yet day, David sware, saying, So do God to me, and more also, if I taste bread, or ought else, till the

Dauid clara adhuc die iurauit Dauid dicens haec faciat mihi sun be down. Deus et haec addat si ante occasum solis gustauero panem uel aliud quicquam 3:36 omnisque populus audiuit et pla- 3:36 And all the people took notice of it, and it pleased

them: as whatsoever the king did pleased all the people.

cuerunt eis cuncta quae fecit rex in conspectu totius populi 3:37 et cognouit omne uulgus et uniuersus Israhel in die illa 3:37 For all the people and all Israel understood that day

that it was not of the king to slay Abner the son of Ner.

quoniam non actum fuisset a rege ut occideretur Abner fi-

lius Ner 3:38 dixit quoque rex ad seruos suos num igno- 3:38 And the king said unto his servants, Know ye not that

there is a prince and a great man fallen this day in Israel?

ratis quoniam princeps et maximus cecidit hodie in Israhel 3:39 ego autem adhuc delicatus et unctus rex porro uiri isti fi- 3:39 And I am this day weak, though anointed king;

and these men the sons of Zeruiah be too hard for me: the LORD shall reward the doer of evil according to his

lii Saruiae duri mihi sunt retribuat Dominus facienti malum wickedness. iuxta malitiam suam

4:1 audiuit autem filius Saul quod cecidisset Abner in He- 4:1 And when Saul’s son heard that Abner was dead in

Hebron, his hands were feeble, and all the Israelites were troubled.

bron et dissolutae sunt manus eius omnisque Israhel perturbatus est 4:2 duo autem uiri principes latronum erant filio 4:2 And Saul’s son had two men that were captains of

bands: the name of the one was Baanah, and the name of the other Rechab, the sons of Rimmon a Beerothite, of the children of Benjamin: (for Beeroth also was reckoned to Benjamin.

Saul nomen uni Baana et nomen alteri Rechab filii Remmon Berothitae de filiis Beniamin siquidem et Beroth reputata est in Beniamin 4:3 et fugerunt Berothitae in Gett- 4:3 And the Beerothites fled to Gittaim, and were sojourn-

ers there until this day.)

haim fueruntque ibi aduenae usque in tempus illud 4:4 erat 4:4 And Jonathan, Saul’s son, had a son that was lame of

his feet. He was five years old when the tidings came of Saul and Jonathan out of Jezreel, and his nurse took him up, and fled: and it came to pass, as she made haste to flee, that he fell, and became lame. And his name was Mephibosheth.

autem Ionathan filio Saul filius debilis pedibus quinquennis enim fuit quando uenit nuntius de Saul et Ionathan ex Iezrahel tollens itaque eum nutrix sua fugit cumque festinaret

Biblia Sacra Vulgata

ut fugeret cecidit et claudus effectus est habuitque uocabu-

4:5 And the sons of Rimmon the Beerothite, Rechab and lum Mifiboseth 4:5 uenientes igitur filii Remmon Berothi-

Baanah, went, and came about the heat of the day to the house of Ishbosheth, who lay on a bed at noon.

tae Rechab et Baana ingressi sunt feruente die domum His-

4:6 And they came thither into the midst of the house, as boseth qui dormiebat super stratum suum meridie 4:6 in-

though they would have fetched wheat; and they smote him under the fifth rib: and Rechab and Baanah his brother

escaped. gressi sunt autem domum adsumentes spicas tritici et per-

cusserunt eum in inguine Rechab et Baana frater eius et fu-

4:7 For when they came into the house, he lay on his bed gerunt 4:7 cum autem ingressi fuissent domum ille dormie-

in his bedchamber, and they smote him, and slew him, and beheaded him, and took his head, and gat them away

through the plain all night. bat super lectulum suum in conclaui et percutientes interfe-

cerunt eum sublatoque capite eius abierunt per uiam deserti

4:8 And they brought the head of Ishbosheth unto David tota nocte 4:8 et adtulerunt caput Hisboseth ad Dauid in

to Hebron, and said to the king, Behold the head of Ishbosheth the son of Saul thine enemy, which sought thy life; and the LORD hath avenged my lord the king this day of Saul, and of his seed.

Hebron dixeruntque ad regem ecce caput Hisboseth filii Saul inimici tui qui quaerebat animam tuam et dedit Dominus do-

4:9 And David answered Rechab and Baanah his brother, mino meo regi ultiones hodie de Saul et de semine eius 4:9

the sons of Rimmon the Beerothite, and said unto them, As the LORD liveth, who hath redeemed my soul out of

all adversity, respondens autem Dauid Rechab et Baana fratri eius filiis

Remmon Berothei dixit ad eos uiuit Dominus qui eruit ani-

4:10 When one told me, saying, Behold, Saul is dead, mam meam de omni angustia 4:10 quoniam eum qui adthinking to have brought good tidings, I took hold of him, and slew him in Ziklag, who thought that I would have

given him a reward for his tidings: nuntiauerat mihi et dixerat mortuus est Saul qui putabat se

prospera nuntiare tenui et occidi in Siceleg cui oportebat me

4:11 How much more, when wicked men have slain a righ- dare mercedem pro nuntio 4:11 quanto magis nunc cum ho-

teous person in his own house upon his bed? shall I not therefore now require his blood of your hand, and take

you away from the earth? mines impii interfecerint uirum innoxium in domo sua su-

per lectulum suum non quaeram sanguinem eius de manu

4:12 And David commanded his young men, and they uestra et auferam uos de terra 4:12 praecepit itaque Dauid

slew them, and cut off their hands and their feet, and hanged them up over the pool in Hebron. But they took the head of Ishbosheth, and buried it in the sepulchre of Abner in Hebron.

pueris et interfecerunt eos praecidentesque manus et pedes eorum suspenderunt eos super piscinam in Hebron caput autem Hisboseth tulerunt et sepelierunt in sepulchro Abner in Hebron

5:1 Then came all the tribes of Israel to David unto He- 5:1 et uenerunt uniuersae tribus Israhel ad Dauid in Heb-

bron, and spake, saying, Behold, we are thy bone and thy flesh.

ron dicentes ecce nos os tuum et caro tua sumus 5:2 sed et

5:2 Also in time past, when Saul was king over us, thou wast he that leddest out and broughtest in Israel: and the LORD said to thee, Thou shalt feed my people Israel, and thou shalt be a captain over Israel.

heri et nudius tertius cum esset Saul rex super nos tu eras educens et reducens Israhel dixit autem Dominus ad te tu pasces populum meum Israhel et tu eris dux super Israhel

5:3 So all the elders of Israel came to the king to Hebron; 5:3 uenerunt quoque et senes de Israhel ad regem in Hebron

and king David made a league with them in Hebron before the LORD: and they anointed David king over Israel.

et percussit cum eis rex Dauid foedus in Hebron coram Do-

Liber II Samuelis

mino unxeruntque Dauid in regem super Israhel 5:4 filius 5:4 David was thirty years old when he began to reign, and

he reigned forty years.

triginta annorum erat Dauid cum regnare coepisset et quadraginta annis regnauit 5:5 in Hebron regnauit super Iudam 5:5 In Hebron he reigned over Judah seven years and six

months: and in Jerusalem he reigned thirty and three years over all Israel and Judah.

septem annis et sex mensibus in Hierusalem autem regnauit triginta tribus annis super omnem Israhel et Iudam 5:6 5:6 And the king and his men went to Jerusalem unto the

Jebusites, the inhabitants of the land: which spake unto David, saying, Except thou take away the blind and the lame, thou shalt not come in hither: thinking, David cannot come in hither.

et abiit rex et omnes uiri qui erant cum eo in Hierusalem ad Iebuseum habitatorem terrae dictumque est ad Dauid ab eis non ingredieris huc nisi abstuleris caecos et claudos di-

centes non ingredietur Dauid huc 5:7 cepit autem Dauid 5:7 Nevertheless David took the strong hold of Zion: the

same is the city of David.

arcem Sion haec est ciuitas Dauid 5:8 proposuerat enim in 5:8 And David said on that day, Whosoever getteth up to

the gutter, and smiteth the Jebusites, and the lame and the blind that are hated of David’s soul, he shall be chief and captain. Wherefore they said, The blind and the lame shall not come into the house.

die illa praemium qui percussisset Iebuseum et tetigisset domatum fistulas et claudos et caecos odientes animam Dauid idcirco dicitur in prouerbio caecus et claudus non intrabunt templum 5:9 habitauit autem Dauid in arce et uocauit eam 5:9 So David dwelt in the fort, and called it the city of

David. And David built round about from Millo and inward.

ciuitatem Dauid et aedificauit per gyrum a Mello et intrinsecus 5:10 et ingrediebatur proficiens atque succrescens et 5:10 And David went on, and grew great, and the LORD

God of hosts was with him.

Dominus Deus exercituum erat cum eo 5:11 misit quoque 5:11 And Hiram king of Tyre sent messengers to David,

and cedar trees, and carpenters, and masons: and they built David an house.

Hiram rex Tyri nuntios ad Dauid et ligna cedrina et artifices lignorum artificesque lapidum ad parietes et aedificauerunt

domum Dauid 5:12 et cognouit Dauid quoniam confirmas- 5:12 And David perceived that the LORD had established him king over Israel, and that he had exalted his kingdom for his people Israel’s sake.

set eum Dominus regem super Israhel et quoniam exaltasset regnum eius super populum suum Israhel 5:13 accepit ergo 5:13 And David took him more concubines and wives out

of Jerusalem, after he was come from Hebron: and there were yet sons and daughters born to David.

adhuc concubinas et uxores de Hierusalem postquam uenerat de Hebron natique sunt Dauid et alii filii et filiae 5:14 et 5:14 And these be the names of those that were born unto

him in Jerusalem; Shammuah, and Shobab, and Nathan, and Solomon,

haec nomina eorum qui nati sunt ei in Hierusalem Samua et Sobab et Nathan et Salomon 5:15 et Ibaar et Helisua et Ne- 5:15 Ibhar also, and Elishua, and Nepheg, and Japhia, pheg 5:16 et Iafia et Helisama et Helida et Helifeleth 5:17 5:16 And Elishama, and Eliada, and Eliphalet.

5:17 But when the Philistines heard that they had anointed David king over Israel, all the Philistines came up to seek David; and David heard of it, and went down to the hold.

audierunt uero Philisthim quod unxissent Dauid regem super Israhel et ascenderunt uniuersi ut quaererent Dauid quod cum audisset Dauid descendit in praesidium 5:18 Philist- 5:18 The Philistines also came and spread themselves in

the valley of Rephaim.

him autem uenientes diffusi sunt in ualle Raphaim 5:19 et 5:19 And David enquired of the LORD, saying, Shall I

go up to the Philistines? wilt thou deliver them into mine hand? And the LORD said unto David, Go up: for I will

consuluit Dauid Dominum dicens si ascendam ad Philist- doubtless deliver the Philistines into thine hand. him et si dabis eos in manu mea et dixit Dominus ad Dauid

Biblia Sacra Vulgata

5:20 And David came to Baalperazim, and David smote ascende quia tradens dabo Philisthim in manu tua 5:20 ue-

them there, and said, The LORD hath broken forth upon mine enemies before me, as the breach of waters. There-

fore he called the name of that place Baalperazim. nit ergo Dauid in Baalpharasim et percussit eos ibi et dixit diuisit Dominus inimicos meos coram me sicut diuiduntur aquae propterea uocatum est nomen loci illius Baalpharasim

5:21 And there they left their images, and David and his 5:21 et reliquerunt ibi sculptilia sua quae tulit Dauid et uiri

men burned them.

5:22 And the Philistines came up yet again, and spread eius 5:22 et addiderunt adhuc Philisthim ut ascenderent et

themselves in the valley of Rephaim.

5:23 And when David enquired of the LORD, he said, diffusi sunt in ualle Raphaim 5:23 consuluit autem Dauid

Thou shalt not go up; but fetch a compass behind them, and come upon them over against the mulberry trees.

Dominum qui respondit non ascendas sed gyra post tergum

5:24 And let it be, when thou hearest the sound of a going eorum et uenies ad eos ex aduerso pirorum 5:24 et cum auin the tops of the mulberry trees, that then thou shalt bestir thyself: for then shall the LORD go out before thee, to

smite the host of the Philistines. dieris sonitum gradientis in cacumine pirorum tunc inibis

proelium quia tunc egredietur Dominus ante faciem tuam ut

5:25 And David did so, as the LORD had commanded percutiat castra Philisthim 5:25 fecit itaque Dauid sicut ei

him; and smote the Philistines from Geba until thou come to Gazer.

praeceperat Dominus et percussit Philisthim de Gabee usque dum uenias Gezer

6:1 Again, David gathered together all the chosen men of 6:1 congregauit autem rursum Dauid omnes electos ex Is-Israel, thirty thousand.

6:2 And David arose, and went with all the people that rahel triginta milia 6:2 surrexitque et abiit et uniuersus po-

were with him from Baale of Judah, to bring up from thence the ark of God, whose name is called by the name of the LORD of hosts that dwelleth between the cherubims.

pulus qui erat cum eo de uiris Iuda ut adducerent arcam Dei super quam inuocatum est nomen Domini exercituum seden-

6:3 And they set the ark of God upon a new cart, and tis in cherubin super eam 6:3 et inposuerunt arcam Domini

brought it out of the house of Abinadab that was in Gibeah: and Uzzah and Ahio, the sons of Abinadab, drave

the new cart. super plaustrum nouum tuleruntque eam de domo Abinadab

qui erat in Gabaa Oza autem et Haio filii Abinadab mina-

6:4 And they brought it out of the house of Abinadab bant plaustrum nouum 6:4 cumque tulissent eam de domo

which was at Gibeah, accompanying the ark of God: and Ahio went before the ark.

Abinadab qui erat in Gabaa custodiens arcam Dei Haio prae-

6:5 And David and all the house of Israel played before cedebat arcam 6:5 Dauid autem et omnis Israhel ludebant

the LORD on all manner of instruments made of fir wood, even on harps, and on psalteries, and on timbrels, and on

cornets, and on cymbals. coram Domino in omnibus lignis fabrefactis et citharis et ly-6:6 And when they came to Nachon’s threshingfloor, Uz- ris et tympanis et sistris et cymbalis 6:6 postquam autem

zah put forth his hand to the ark of God, and took hold of it; for the oxen shook it.

uenerunt ad aream Nachon extendit manum Oza ad arcam

6:7 And the anger of the LORD was kindled against Uz- Dei et tenuit eam quoniam calcitrabant boues 6:7 iratus-

zah; and God smote him there for his error; and there he died by the ark of God.

que est indignatione Dominus contra Ozam et percussit eum

6:8 And David was displeased, because the LORD had super temeritate qui mortuus est ibi iuxta arcam Dei 6:8

made a breach upon Uzzah: and he called the name of the place Perezuzzah to this day.

contristatus autem est Dauid eo quod percussisset Dominus Ozam et uocatum est nomen loci illius Percussio Oza us-

6:9 And David was afraid of the LORD that day, and said, que in diem hanc 6:9 et extimuit Dauid Dominum in die How shall the ark of the LORD come to me?

Liber II Samuelis

illa dicens quomodo ingredietur ad me arca Domini 6:10 et 6:10 So David would not remove the ark of the LORD

unto him into the city of David: but David carried it aside into the house of Obededom the Gittite.

noluit deuertere ad se arcam Domini in ciuitate Dauid sed deuertit eam in domo Obededom Getthei 6:11 et habitauit 6:11 And the ark of the LORD continued in the house

of Obededom the Gittite three months: and the LORD blessed Obededom, and all his household.

arca Domini in domo Obededom Getthei tribus mensibus et benedixit Dominus Obededom et omnem domum eius 6:12 6:12 And it was told king David, saying, The LORD hath

blessed the house of Obededom, and all that pertaineth unto him, because of the ark of God. So David went and brought up the ark of God from the house of Obededom into the city of David with gladness.

nuntiatumque est regi Dauid benedixit Dominus Obededom et omnia eius propter arcam Dei abiit ergo Dauid et adduxit arcam Dei de domo Obededom in ciuitatem Dauid cum gaudio 6:13 cumque transcendissent qui portabant arcam Do- 6:13 And it was so, that when they that bare the ark of the

LORD had gone six paces, he sacrificed oxen and fatlings.

mini sex passus immolabat bouem et arietem 6:14 et Da- 6:14 And David danced before the LORD with all his

might; and David was girded with a linen ephod.

uid saltabat totis uiribus ante Dominum porro Dauid erat accinctus ephod lineo 6:15 et Dauid et omnis domus Israhel 6:15 So David and all the house of Israel brought up the

ark of the LORD with shouting, and with the sound of the trumpet.

ducebant arcam testamenti Domini in iubilo et in clangore bucinae 6:16 cumque intrasset arca Domini ciuitatem Da- 6:16 And as the ark of the LORD came into the city

of David, Michal Saul’s daughter looked through a window, and saw king David leaping and dancing before the

uid Michol filia Saul prospiciens per fenestram uidit regem LORD; and she despised him in her heart. Dauid subsilientem atque saltantem coram Domino et despexit eum in corde suo 6:17 et introduxerunt arcam Do- 6:17 And they brought in the ark of the LORD, and set it

in his place, in the midst of the tabernacle that David had pitched for it: and David offered burnt offerings and peace

mini et posuerunt eam in loco suo in medio tabernaculi quod offerings before the LORD. tetenderat ei Dauid et obtulit Dauid holocausta coram Domino et pacifica 6:18 cumque conplesset offerens holocau- 6:18 And as soon as David had made an end of offering

burnt offerings and peace offerings, he blessed the people in the name of the LORD of hosts.

stum et pacifica benedixit populo in nomine Domini exercituum 6:19 et partitus est multitudini uniuersae Israhel tam 6:19 And he dealt among all the people, even among the

whole multitude of Israel, as well to the women as men, to every one a cake of bread, and a good piece of flesh, and a flagon of wine. So all the people departed every one to his house.

uiro quam mulieri singulis collyridam panis unam et assaturam bubulae carnis unam et similam frixam oleo et abiit omnis populus unusquisque in domum suam 6:20 reuersusque 6:20 Then David returned to bless his household. And

Michal the daughter of Saul came out to meet David, and said, How glorious was the king of Israel to day, who uncovered himself to day in the eyes of the handmaids of his servants, as one of the vain fellows shamelessly uncovereth himself!

est et Dauid ut benediceret domui suae et egressa Michol filia Saul in occursum Dauid ait quam gloriosus fuit hodie rex Israhel discoperiens se ante ancillas seruorum suorum et nudatus est quasi si nudetur unus de scurris 6:21 dixitque 6:21 And David said unto Michal, It was before the

LORD, which chose me before thy father, and before all his house, to appoint me ruler over the people of the

Dauid ad Michol ante Dominum qui elegit me potius quam LORD, over Israel: therefore will I play before the LORD.

patrem tuum et quam omnem domum eius et praecepit mihi ut essem dux super populum Domini Israhel 6:22 et ludam 6:22 And I will yet be more vile than thus, and will be

base in mine own sight: and of the maidservants which thou hast spoken of, of them shall I be had in honour.

et uilior fiam plus quam factus sum et ero humilis in oculis

Biblia Sacra Vulgata

meis et cum ancillis de quibus locuta es gloriosior apparebo

6:23 Therefore Michal the daughter of Saul had no child 6:23 igitur Michol filiae Saul non est natus filius usque ad

unto the day of her death.

diem mortis suae

7:1 And it came to pass, when the king sat in his house, 7:1 factum est autem cum sedisset rex in domo sua et Do-

and the LORD had given him rest round about from all his enemies;

minus dedisset ei requiem undique ab uniuersis inimicis suis

7:2 That the king said unto Nathan the prophet, See now, 7:2 dixit ad Nathan prophetam uidesne quod ego habitem in

I dwell in an house of cedar, but the ark of God dwelleth within curtains.

domo cedrina et arca Dei posita sit in medio pellium 7:3

7:3 And Nathan said to the king, Go, do all that is in thine heart; for the LORD is with thee.

dixitque Nathan ad regem omne quod est in corde tuo uade

7:4 And it came to pass that night, that the word of the fac quia Dominus tecum est 7:4 factum est autem in nocte

LORD came unto Nathan, saying,

7:5 Go and tell my servant David, Thus saith the LORD, illa et ecce sermo Domini ad Nathan dicens 7:5 uade et lo-

Shalt thou build me an house for me to dwell in?

quere ad seruum meum Dauid haec dicit Dominus numquid

7:6 Whereas I have not dwelt in any house since the time tu aedificabis mihi domum ad habitandum 7:6 neque enim that I brought up the children of Israel out of Egypt, even to this day, but have walked in a tent and in a tabernacle.

habitaui in domo ex die qua eduxi filios Israhel de terra Aegypti usque in diem hanc sed ambulans ambulabam in taber-

7:7 In all the places wherein I have walked with all the naculo et in tentorio 7:7 per cuncta loca quae transiui cum

children of Israel spake I a word with any of the tribes of Israel, whom I commanded to feed my people Israel,

saying, Why build ye not me an house of cedar? omnibus filiis Israhel numquid loquens locutus sum ad unam de tribubus Israhel cui praecepi ut pasceret populum meum Israhel dicens quare non aedificastis mihi domum cedrinam

7:8 Now therefore so shalt thou say unto my servant 7:8 et nunc haec dices seruo meo Dauid haec dicit Dominus

David, Thus saith the LORD of hosts, I took thee from the sheepcote, from following the sheep, to be ruler over

my people, over Israel: exercituum ego tuli te de pascuis sequentem greges ut es-

7:9 And I was with thee whithersoever thou wentest, and ses dux super populum meum Israhel 7:9 et fui tecum in have cut off all thine enemies out of thy sight, and have made thee a great name, like unto the name of the great

men that are in the earth. omnibus ubicumque ambulasti et interfeci uniuersos inimi-

cos tuos a facie tua fecique tibi nomen grande iuxta nomen

7:10 Moreover I will appoint a place for my people Is- magnorum qui sunt in terra 7:10 et ponam locum populo rael, and will plant them, that they may dwell in a place of their own, and move no more; neither shall the children of

wickedness afflict them any more, as beforetime, meo Israhel et plantabo eum et habitabit sub eo et non tur-

babitur amplius nec addent filii iniquitatis ut adfligant eum

7:11 And as since the time that I commanded judges to be sicut prius 7:11 ex die qua constitui iudices super populum

over my people Israel, and have caused thee to rest from all thine enemies. Also the LORD telleth thee that he will

make thee an house. meum Israhel et requiem dabo tibi ab omnibus inimicis tuis

praedicitque tibi Dominus quod domum faciat tibi Dominus

7:12 And when thy days be fulfilled, and thou shalt sleep 7:12 cumque conpleti fuerint dies tui et dormieris cum pat-

with thy fathers, I will set up thy seed after thee, which shall proceed out of thy bowels, and I will establish his

kingdom. ribus tuis suscitabo semen tuum post te quod egredietur de 7:13 He shall build an house for my name, and I will stab- utero tuo et firmabo regnum eius 7:13 ipse aedificabit do-

lish the throne of his kingdom for ever.

mum nomini meo et stabiliam thronum regni eius usque in

Liber II Samuelis

sempiternum 7:14 ego ero ei in patrem et ipse erit mihi in 7:14 I will be his father, and he shall be my son. If he

commit iniquity, I will chasten him with the rod of men, and with the stripes of the children of men:

filium qui si inique aliquid gesserit arguam eum in uirga uirorum et in plagis filiorum hominum 7:15 misericordiam 7:15 But my mercy shall not depart away from him, as I

took it from Saul, whom I put away before thee.

autem meam non auferam ab eo sicut abstuli a Saul quem

amoui a facie tua 7:16 et fidelis erit domus tua et regnum 7:16 And thine house and thy kingdom shall be established for ever before thee: thy throne shall be established for ever.

tuum usque in aeternum ante faciem tuam et thronus tuus

erit firmus iugiter 7:17 secundum omnia uerba haec et iuxta 7:17 According to all these words, and according to all this

vision, so did Nathan speak unto David.

uniuersam uisionem istam sic locutus est Nathan ad Dauid 7:18 ingressus est autem rex Dauid et sedit coram Domino et 7:18 Then went king David in, and sat before the LORD,

and he said, Who am I, O Lord GOD? and what is my house, that thou hast brought me hitherto?

dixit quis ego sum Domine Deus et quae domus mea quia adduxisti me hucusque 7:19 sed et hoc parum uisum est in 7:19 And this was yet a small thing in thy sight, O Lord

GOD; but thou hast spoken also of thy servant’s house for a great while to come. And is this the manner of man, O

conspectu tuo Domine Deus nisi loquereris etiam de domo Lord GOD? serui tui in longinquum ista est enim lex Adam Domine Deus 7:20 quid ergo addere poterit adhuc Dauid ut loquatur ad te tu 7:20 And what can David say more unto thee? for thou,

Lord GOD, knowest thy servant.

enim scis seruum tuum Domine Deus 7:21 propter uerbum 7:21 For thy word’s sake, and according to thine own

heart, hast thou done all these great things, to make thy servant know them.

tuum et secundum cor tuum fecisti omnia magnalia haec ita ut notum faceres seruo tuo 7:22 idcirco magnificatus es Do- 7:22 Wherefore thou art great, O LORD God: for there is

none like thee, neither is there any God beside thee, according to all that we have heard with our ears.

mine Deus quia non est similis tui neque est deus extra te in omnibus quae audiuimus auribus nostris 7:23 quae est au- 7:23 And what one nation in the earth is like thy people,

even like Israel, whom God went to redeem for a people to himself, and to make him a name, and to do for you great things and terrible, for thy land, before thy people, which thou redeemedst to thee from Egypt, from the nations and their gods?

tem ut populus tuus Israhel gens in terra propter quam iuit Deus ut rediieret eam sibi in populum et poneret sibi nomen faceretque eis magnalia et horribilia super terram a facie populi tui quem redeeisti tibi ex Aegypto gentem et deum eius 7:24 et firmasti tibi populum tuum Israhel in populum sempi- 7:24 For thou hast confirmed to thyself thy people Israel

to be a people unto thee for ever: and thou, LORD, art become their God.

ternum et tu Domine factus es eis in Deum 7:25 nunc ergo

7:25 And now, O LORD God, the word that thou hast spoken concerning thy servant, and concerning his house, establish it for ever, and do as thou hast said.

Domine Deus uerbum quod locutus es super seruum tuum et super domum eius suscita in sempiternum et fac sicut lo-

cutus es 7:26 et magnificetur nomen tuum usque in sem- 7:26 And let thy name be magnified for ever, saying, The LORD of hosts is the God over Israel: and let the house of thy servant David be established before thee.

piternum atque dicatur Dominus exercituum Deus super Israhel et domus serui tui Dauid erit stabilita coram Domino 7:27 quia tu Domine exercituum Deus Israhel reuelasti au- 7:27 For thou, O LORD of hosts, God of Israel, hast re-

vealed to thy servant, saying, I will build thee an house: therefore hath thy servant found in his heart to pray this

rem serui tui dicens domum aedificabo tibi propterea inuenit prayer unto thee. seruus tuus cor suum ut oraret te oratione hac 7:28 nunc 7:28 And now, O Lord GOD, thou art that God, and thy

words be true, and thou hast promised this goodness unto thy servant:

Biblia Sacra Vulgata

ergo Domine Deus tu es Deus et uerba tua erunt uera locu-

7:29 Therefore now let it please thee to bless the house of tus es enim ad seruum tuum bona haec 7:29 incipe igitur et thy servant, that it may continue for ever before thee: for thou, O Lord GOD, hast spoken it: and with thy blessing

let the house of thy servant be blessed for ever. benedic domui serui tui ut sit in sempiternum coram te quia tu Domine Deus locutus es et benedictione tua benedicetur domus serui tui in sempiternum

8:1 And after this it came to pass that David smote 8:1 factum est autem post haec percussit Dauid Philisthim

the Philistines, and subdued them: and David took Methegammah out of the hand of the Philistines.

et humiliauit eos et tulit Dauid frenum tributi de manu Phi-

8:2 And he smote Moab, and measured them with a line, listhim 8:2 et percussit Moab et mensus est eos funiculo

casting them down to the ground; even with two lines measured he to put to death, and with one full line to keep alive. And so the Moabites became David’s servants, and brought gifts.

coaequans terrae mensus est autem duos funiculos unum ad occidendum et unum ad uiuificandum factusque est Moab

8:3 David smote also Hadadezer, the son of Rehob, king Dauid seruiens sub tributo 8:3 et percussit Dauid Adadezer

of Zobah, as he went to recover his border at the river Euphrates.

filium Roob regem Soba quando profectus est ut dominare-

8:4 And David took from him a thousand chariots, and tur super flumen Eufraten 8:4 et captis Dauid ex parte eius

seven hundred horsemen, and twenty thousand footmen: and David houghed all the chariot horses, but reserved of

them for an hundred chariots. mille septingentis equitibus et uiginti milibus peditum sub-

neruauit omnes iugales curruum dereliquit autem ex eis cen-

8:5 And when the Syrians of Damascus came to succour tum currus 8:5 uenit quoque Syria Damasci ut praesidium

Hadadezer king of Zobah, David slew of the Syrians two and twenty thousand men.

ferret Adadezer regi Soba et percussit Dauid de Syria uiginti

8:6 Then David put garrisons in Syria of Damascus: and duo milia uirorum 8:6 et posuit Dauid praesidium in Syria

the Syrians became servants to David, and brought gifts. And the LORD preserved David whithersoever he went.

Damasci factaque est Syria Dauid seruiens sub tributo seruauit Dominus Dauid in omnibus ad quaecumque profectus est

8:7 And David took the shields of gold that were on the 8:7 et tulit Dauid arma aurea quae habebant serui Adadezer et

servants of Hadadezer, and brought them to Jerusalem.

8:8 And from Betah, and from Berothai, cities of detulit ea in Hierusalem 8:8 et de Bete et de Beroth ciuitati-

Hadadezer, king David took exceeding much brass.

8:9 When Toi king of Hamath heard that David had smit- bus Adadezer tulit rex Dauid aes multum nimis 8:9 audiuit

ten all the host of Hadadezer,

autem Thou rex Emath quod percussisset Dauid omne robur

8:10 Then Toi sent Joram his son unto king David, to Adadezer 8:10 et misit Thou Ioram filium suum ad regem

salute him, and to bless him, because he had fought against Hadadezer, and smitten him: for Hadadezer had wars with Toi. And Joram brought with him vessels of silver, and vessels of gold, and vessels of brass:

Dauid ut salutaret eum congratulans et gratias ageret eo quod expugnasset Adadezer et percussisset eum hostis quippe erat Thou Adadezer et in manu eius erant uasa argentea et uasa

8:11 Which also king David did dedicate unto the LORD, aurea et uasa aerea 8:11 quae et ipsa sanctificauit rex Dauid

with the silver and gold that he had dedicated of all nations which he subdued;

Domino cum argento et auro quae sanctificauerat de uniuer-

8:12 Of Syria, and of Moab, and of the children of Am- sis gentibus quas subegerat 8:12 de Syria et Moab et filiis

mon, and of the Philistines, and of Amalek, and of the spoil of Hadadezer, son of Rehob, king of Zobah.

Ammon et Philisthim et Amalech et de manubiis Adadezer

8:13 And David gat him a name when he returned from filii Roob regis Soba 8:13 fecit quoque sibi Dauid nomen

smiting of the Syrians in the valley of salt, being eighteen thousand men.

Liber II Samuelis

cum reuerteretur capta Syria in ualle Salinarum caesis duodecim milibus 8:14 et posuit in Idumea custodes statuitque 8:14 And he put garrisons in Edom; throughout all Edom

put he garrisons, and all they of Edom became David’s servants. And the LORD preserved David whithersoever

praesidium et facta est uniuersa Idumea seruiens Dauid et he went. seruauit Dominus Dauid in omnibus ad quaecumque profectus est 8:15 et regnauit Dauid super omnem Israhel faciebat 8:15 And David reigned over all Israel; and David exe-

cuted judgment and justice unto all his people.

quoque Dauid iudicium et iustitiam omni populo suo 8:16 8:16 And Joab the son of Zeruiah was over the host; and

Jehoshaphat the son of Ahilud was recorder;

Ioab autem filius Saruiae erat super exercitum porro Iosaphat filius Ahilud erat a commentariis 8:17 et Sadoc filius 8:17 And Zadok the son of Ahitub, and Ahimelech the son

of Abiathar, were the priests; and Seraiah was the scribe;

Achitob et Ahimelech filius Abiathar sacerdotes et Saraias scriba 8:18 Banaias autem filius Ioiada super Cherethi et 8:18 And Benaiah the son of Jehoiada was over both the

Cherethites and the Pelethites; and David’s sons were chief rulers.

Felethi filii autem Dauid sacerdotes erant

9:1 et dixit Dauid putasne est aliquis qui remanserit de 9:1 And David said, Is there yet any that is left of the house

of Saul, that I may shew him kindness for Jonathan’s sake?

domo Saul ut faciam cum eo misericordiam propter Ionathan 9:2 erat autem de domo Saul seruus nomine Siba quem 9:2 And there was of the house of Saul a servant whose

name was Ziba. And when they had called him unto David, the king said unto him, Art thou Ziba? And he

cum uocasset rex ad se dixit ei tune es Siba et ille respon- said, Thy servant is he. dit ego sum seruus tuus 9:3 et ait rex num superest aliquis 9:3 And the king said, Is there not yet any of the house of

Saul, that I may shew the kindness of God unto him? And Ziba said unto the king, Jonathan hath yet a son, which is

de domo Saul ut faciam cum eo misericordiam Dei dixitque lame on his feet. Siba regi superest filius Ionathan debilis pedibus 9:4 ubi 9:4 And the king said unto him, Where is he? And Ziba

said unto the king, Behold, he is in the house of Machir, the son of Ammiel, in Lodebar.

inquit est et Siba ad regem ecce ait in domo est Machir filii Amihel in Lodabar 9:5 misit ergo rex Dauid et tulit eum de 9:5 Then king David sent, and fetched him out of the house

of Machir, the son of Ammiel, from Lodebar.

domo Machir filii Amihel de Lodabar 9:6 cum autem ue- 9:6 Now when Mephibosheth, the son of Jonathan, the son of Saul, was come unto David, he fell on his face, and did reverence. And David said, Mephibosheth. And he

nisset Mifiboseth filius Ionathan filii Saul ad Dauid corruit answered, Behold thy servant! in faciem suam et adorauit dixitque Dauid Mifiboseth qui respondit adsum seruus tuus 9:7 et ait ei Dauid ne timeas 9:7 And David said unto him, Fear not: for I will surely

shew thee kindness for Jonathan thy father’s sake, and will restore thee all the land of Saul thy father; and thou shalt

quia faciens faciam in te misericordiam propter Ionathan pa- eat bread at my table continually. trem tuum et restituam tibi omnes agros Saul patris tui et tu

comedes panem in mensa mea semper 9:8 qui adorans eum 9:8 And he bowed himself, and said, What is thy servant,

that thou shouldest look upon such a dead dog as I am?

dixit quis ego sum seruus tuus quoniam respexisti super ca-

nem mortuum similem mei 9:9 uocauit itaque rex Sibam 9:9 Then the king called to Ziba, Saul’s servant, and said unto him, I have given unto thy master’s son all that pertained to Saul and to all his house.

puerum Saul et dixit ei omnia quaecumque fuerunt Saul et

uniuersam domum eius dedi filio domini tui 9:10 operare 9:10 Thou therefore, and thy sons, and thy servants, shall till the land for him, and thou shalt bring in the fruits, that thy master’s son may have food to eat: but Mephibosheth thy master’s son shall eat bread alway at my table. Now Ziba had fifteen sons and twenty servants.

igitur ei terram tu et filii tui et serui tui et inferes filio domini tui cibos ut alatur Mifiboseth autem filius domini tui come-

Biblia Sacra Vulgata

det semper panem super mensam meam erant autem Sibae

9:11 Then said Ziba unto the king, According to all that quindecim filii et uiginti serui 9:11 dixitque Siba ad regem

my lord the king hath commanded his servant, so shall thy servant do. As for Mephibosheth, said the king, he shall

eat at my table, as one of the king’s sons. sicut iussisti domine mi rex seruo tuo sic faciet seruus tuus

et Mifiboseth comedet super mensam tuam quasi unus de fi-

9:12 And Mephibosheth had a young son, whose name liis regis 9:12 habebat autem Mifiboseth filium paruulum

was Micha. And all that dwelt in the house of Ziba were servants unto Mephibosheth.

nomine Micha omnis uero cognatio domus Siba seruiebat

9:13 So Mephibosheth dwelt in Jerusalem: for he did eat Mifiboseth 9:13 porro Mifiboseth habitabat in Hierusalem

continually at the king’s table; and was lame on both his feet.

quia de mensa regis iugiter uescebatur et erat claudus utroque pede

10:1 And it came to pass after this, that the king of the 10:1 factum est autem post haec ut moreretur rex filiorum

children of Ammon died, and Hanun his son reigned in his stead.

Ammon et regnaret Anon filius eius pro eo 10:2 dixitque

10:2 Then said David, I will shew kindness unto Hanun the son of Nahash, as his father shewed kindness unto me. And David sent to comfort him by the hand of his servants for his father. And David’s servants came into the land of the children of Ammon.

Dauid faciam misericordiam cum Anon filio Naas sicut fecit pater eius mecum misericordiam misit ergo Dauid consolans eum per seruos suos super patris interitu cum autem uenis-

10:3 And the princes of the children of Ammon said unto sent serui Dauid in terram filiorum Ammon 10:3 dixerunt

Hanun their lord, Thinkest thou that David doth honour thy father, that he hath sent comforters unto thee? hath not David rather sent his servants unto thee, to search the city, and to spy it out, and to overthrow it?

principes filiorum Ammon ad Anon dominum suum putas quod propter honorem patris tui Dauid miserit ad te consolatores et non ideo ut inuestigaret et exploraret ciuitatem et

10:4 Wherefore Hanun took David’s servants, and shaved euerteret eam misit Dauid seruos suos ad te 10:4 tulit itaque

off the one half of their beards, and cut off their garments in the middle, even to their buttocks, and sent them away.

Anon seruos Dauid rasitque dimidiam partem barbae eorum et praecidit uestes eorum medias usque ad nates et dimisit

10:5 When they told it unto David, he sent to meet them, eos 10:5 quod cum nuntiatum esset Dauid misit in occur-

because the men were greatly ashamed: and the king said, Tarry at Jericho until your beards be grown, and then re-

turn. sum eorum erant enim uiri confusi turpiter ualde et manda-

uit eis Dauid manete Hiericho donec crescat barba uestra et

10:6 And when the children of Ammon saw that they stank tunc reuertimini 10:6 uidentes autem filii Ammon quod in-

before David, the children of Ammon sent and hired the Syrians of Bethrehob and the Syrians of Zoba, twenty thousand footmen, and of king Maacah a thousand men, and of Ishtob twelve thousand men.

iuriam fecissent Dauid miserunt et conduxerunt mercede Syrum Roob et Syrum Soba uiginti milia peditum et a rege Ma-

10:7 And when David heard of it, he sent Joab, and all the acha mille uiros et ab Histob duodecim milia uirorum 10:7

host of the mighty men.

quod cum audisset Dauid misit Ioab et omnem exercitum

10:8 And the children of Ammon came out, and put the bellatorum 10:8 egressi sunt ergo filii Ammon et direxe-

battle in array at the entering in of the gate: and the Syrians of Zoba, and of Rehob, and Ishtob, and Maacah, were by

themselves in the field. runt aciem ante ipsum introitum portae Syrus autem Soba

10:9 When Joab saw that the front of the battle was against et Roob et Histob et Maacha seorsum erant in campo 10:9 him before and behind, he chose of all the choice men of Israel, and put them in array against the Syrians:

uidens igitur Ioab quod praeparatum esset aduersum se pro-

Liber II Samuelis

elium et ex aduerso et post tergum elegit ex omnibus electis Israhel et instruxit aciem contra Syrum 10:10 reliquam au- 10:10 And the rest of the people he delivered into the hand

of Abishai his brother, that he might put them in array against the children of Ammon.

tem partem populi tradidit Abisai fratri suo qui direxit aciem aduersum filios Ammon 10:11 et ait Ioab si praeualuerint 10:11 And he said, If the Syrians be too strong for me,

then thou shalt help me: but if the children of Ammon be too strong for thee, then I will come and help thee.

aduersum me Syri eris mihi in adiutorium si autem filii Ammon praeualuerint aduersum te auxiliabor tibi 10:12 esto 10:12 Be of good courage, and let us play the men for our

people, and for the cities of our God: and the LORD do that which seemeth him good.

uir fortis et pugnemus pro populo nostro et ciuitate Dei nostri Dominus autem faciet quod bonum est in conspectu suo 10:13 iniit itaque Ioab et populus qui erat cum eo certamen 10:13 And Joab drew nigh, and the people that were with

him, unto the battle against the Syrians: and they fled before him.

contra Syros qui statim fugerunt a facie eius 10:14 filii au-

10:14 And when the children of Ammon saw that the Syrians were fled, then fled they also before Abishai, and entered into the city. So Joab returned from the children of Ammon, and came to Jerusalem.

tem Ammon uidentes quod fugissent Syri fugerunt et ipsi a facie Abisai et ingressi sunt ciuitatem reuersusque est Ioab a filiis Ammon et uenit Hierusalem 10:15 uidentes igitur 10:15 And when the Syrians saw that they were smitten

before Israel, they gathered themselves together.

Syri quoniam corruissent coram Israhel congregati sunt pariter 10:16 misitque Adadezer et eduxit Syros qui erant trans 10:16 And Hadarezer sent, and brought out the Syrians

that were beyond the river: and they came to Helam; and Shobach the captain of the host of Hadarezer went before

Fluuium et adduxit exercitum eorum Sobach autem magis- them. ter militiae Adadezer erat princeps eorum 10:17 quod cum 10:17 And when it was told David, he gathered all Is-

rael together, and passed over Jordan, and came to Helam. And the Syrians set themselves in array against David, and

nuntiatum esset Dauid contraxit omnem Israhelem et trans- fought with him. iuit Iordanem uenitque in Helema et direxerunt aciem Syri ex aduerso Dauid et pugnauerunt contra eum 10:18 fuge- 10:18 And the Syrians fled before Israel; and David slew

the men of seven hundred chariots of the Syrians, and forty thousand horsemen, and smote Shobach the captain

runtque Syri a facie Israhel et occidit Dauid de Syris septin- of their host, who died there. gentos currus et quadraginta milia equitum et Sobach principem militiae percussit qui statim mortuus est 10:19 ui- 10:19 And when all the kings that were servants to

Hadarezer saw that they were smitten before Israel, they made peace with Israel, and served them. So the Syrians

dentes autem uniuersi reges qui erant in praesidio Adadezer feared to help the children of Ammon any more. uictos se ab Israhel fecerunt pacem cum Israhel et seruierunt eis timueruntque Syri auxilium praebere filiis Ammon

11:1 factum est ergo uertente anno eo tempore quo solent 11:1 And it came to pass, after the year was expired, at

the time when kings go forth to battle, that David sent Joab, and his servants with him, and all Israel; and they destroyed the children of Ammon, and besieged Rabbah. But David tarried still at Jerusalem.

reges ad bella procedere misit Dauid Ioab et seruos suos cum eo et uniuersum Israhel et uastauerunt filios Ammon et obsederunt Rabba Dauid autem remansit in Hierusalem 11:2 11:2 And it came to pass in an eveningtide, that David

arose from off his bed, and walked upon the roof of the king’s house: and from the roof he saw a woman washing

dum haec agerentur accidit ut surgeret Dauid de stratu suo herself; and the woman was very beautiful to look upon. post meridiem et deambularet in solario domus regiae uiditque mulierem se lauantem ex aduerso super solarium suum

Biblia Sacra Vulgata

11:3 And David sent and enquired after the woman. And erat autem mulier pulchra ualde 11:3 misit ergo rex et re-

one said, Is not this Bathsheba, the daughter of Eliam, the wife of Uriah the Hittite?

quisiuit quae esset mulier nuntiatumque ei est quod ipsa es-

11:4 And David sent messengers, and took her; and she set Bethsabee filia Heliam uxor Vriae Hetthei 11:4 missis

came in unto him, and he lay with her; for she was purified from her uncleanness: and she returned unto her house.

itaque Dauid nuntiis tulit eam quae cum ingressa esset ad illum dormiuit cum ea statimque sanctificata est ab inmunditia

11:5 And the woman conceived, and sent and told David, sua 11:5 et reuersa est domum suam concepto fetu mittens-

and said, I am with child.

11:6 And David sent to Joab, saying, Send me Uriah the que nuntiauit Dauid et ait concepi 11:6 misit autem Dauid

Hittite. And Joab sent Uriah to David.

ad Ioab dicens mitte ad me Vriam Hettheum misitque Ioab

11:7 And when Uriah was come unto him, David de- Vriam ad Dauid 11:7 et uenit Vrias ad Dauid quaesiuitque

manded of him how Joab did, and how the people did, and how the war prospered.

Dauid quam recte ageret Ioab et populus et quomodo ad-

11:8 And David said to Uriah, Go down to thy house, and ministraretur bellum 11:8 et dixit Dauid ad Vriam uade in

wash thy feet. And Uriah departed out of the king’s house, and there followed him a mess of meat from the king.

domum tuam et laua pedes tuos egressus est Vrias de domo

11:9 But Uriah slept at the door of the king’s house with regis secutusque est eum cibus regius 11:9 dormiuit autem

all the servants of his lord, and went not down to his house.

Vrias ante portam domus regiae cum aliis seruis domini sui

11:10 And when they had told David, saying, Uriah went et non descendit ad domum suam 11:10 nuntiatumque est not down unto his house, David said unto Uriah, Camest thou not from thy journey? why then didst thou not go

down unto thine house? Dauid a dicentibus non iuit Vrias ad domum suam et ait Da-

uid ad Vriam numquid non de uia uenisti quare non descen-

11:11 And Uriah said unto David, The ark, and Israel, and disti ad domum tuam 11:11 et ait Vrias ad Dauid arca et Is-Judah, abide in tents; and my lord Joab, and the servants of my lord, are encamped in the open fields; shall I then go into mine house, to eat and to drink, and to lie with my wife? as thou livest, and as thy soul liveth, I will not do this thing.

rahel et Iuda habitant in papilionibus et dominus meus Ioab et serui domini mei super faciem terrae manent et ego ingrediar domum meam ut comedam et bibam et dormiam cum uxore mea per salutem tuam et per salutem animae tuae quod

11:12 And David said to Uriah, Tarry here to day also, non faciam rem hanc 11:12 ait ergo Dauid ad Vriam mane and to morrow I will let thee depart. So Uriah abode in Jerusalem that day, and the morrow.

hic etiam hodie et cras dimittam te mansit Vrias in Hierusa-

11:13 And when David had called him, he did eat and lem die illa et altera 11:13 et uocauit eum Dauid ut comede-

drink before him; and he made him drunk: and at even he went out to lie on his bed with the servants of his lord,

but went not down to his house. ret coram se et biberet et inebriauit eum qui egressus uespere

dormiuit in stratu suo cum seruis domini sui et in domum

11:14 And it came to pass in the morning, that David wrote suam non descendit 11:14 factum est ergo mane et scripsit

a letter to Joab, and sent it by the hand of Uriah. 11:15 And he wrote in the letter, saying, Set ye Uriah in Dauid epistulam ad Ioab misitque per manum Vriae 11:15 the forefront of the hottest battle, and retire ye from him, that he may be smitten, and die.

scribens in epistula ponite Vriam ex aduerso belli ubi fortissimum proelium est et derelinquite eum ut percussus in-

11:16 And it came to pass, when Joab observed the city, tereat 11:16 igitur cum Ioab obsideret urbem posuit Vriam

that he assigned Uriah unto a place where he knew that valiant men were.

in loco quo sciebat uiros esse fortissimos 11:17 egressique

11:17 And the men of the city went out, and fought with Joab: and there fell some of the people of the servants of David; and Uriah the Hittite died also.

Liber II Samuelis

uiri de ciuitate bellabant aduersum Ioab et ceciderunt de populo seruorum Dauid et mortuus est etiam Vrias Hettheus 11:18 misit itaque Ioab et nuntiauit Dauid omnia uerba proelii 11:18 Then Joab sent and told David all the things con-

cerning the war;

11:19 praecepitque nuntio dicens cum conpleueris uniuersos 11:19 And charged the messenger, saying, When thou hast

made an end of telling the matters of the war unto the king,

sermones belli ad regem 11:20 si eum uideris indignari et 11:20 And if so be that the king’s wrath arise, and he say

unto thee, Wherefore approached ye so nigh unto the city when ye did fight? knew ye not that they would shoot from

dixerit quare accessistis ad murum ut proeliaremini an ig- the wall? norabatis quod multa desuper ex muro tela mittantur 11:21 11:21 Who smote Abimelech the son of Jerubbesheth? did

not a woman cast a piece of a millstone upon him from the wall, that he died in Thebez? why went ye nigh the wall?

quis percussit Abimelech filium Hieroboseth nonne mulier then say thou, Thy servant Uriah the Hittite is dead also. misit super eum fragmen molae de muro et interfecit eum in Thebes quare iuxta murum accessistis dices etiam seruus tuus Vrias Hettheus occubuit 11:22 abiit ergo nuntius et ue- 11:22 So the messenger went, and came and shewed David

all that Joab had sent him for.

nit et narrauit Dauid omnia quae ei praeceperat Ioab 11:23 11:23 And the messenger said unto David, Surely the men

prevailed against us, and came out unto us into the field, and we were upon them even unto the entering of the gate.

et dixit nuntius ad Dauid praeualuerunt aduersum nos uiri et egressi sunt ad nos in agrum nos autem facto impetu persecuti eos sumus usque ad portam ciuitatis 11:24 et direxerunt 11:24 And the shooters shot from off the wall upon thy

servants; and some of the king’s servants be dead, and thy servant Uriah the Hittite is dead also.

iacula sagittarii ad seruos tuos ex muro desuper mortuique sunt de seruis regis quin etiam seruus tuus Vrias Hettheus mortuus est 11:25 et dixit Dauid ad nuntium haec dices Ioab 11:25 Then David said unto the messenger, Thus shalt

thou say unto Joab, Let not this thing displease thee, for the sword devoureth one as well as another: make thy battle more strong against the city, and overthrow it: and encourage thou him.

non te frangat ista res uarius enim euentus est proelii et nunc hunc nunc illum consumit gladius conforta bellatores tuos aduersum urbem ut destruas eam et exhortare eos 11:26 au- 11:26 And when the wife of Uriah heard that Uriah her

husband was dead, she mourned for her husband.

diuit autem uxor Vriae quod mortuus esset Vrias uir suus et planxit eum 11:27 transactoque luctu misit Dauid et in- 11:27 And when the mourning was past, David sent and

fetched her to his house, and she became his wife, and bare him a son. But the thing that David had done displeased

troduxit eam domum suam et facta est ei uxor peperitque ei the LORD. filium et displicuit uerbum hoc quod fecerat Dauid coram Domino

12:1 misit ergo Dominus Nathan ad Dauid qui cum uenis- 12:1 And the LORD sent Nathan unto David. And he

came unto him, and said unto him, There were two men in one city; the one rich, and the other poor.

set ad eum dixit ei duo uiri erant in ciuitate una unus diues et alter pauper 12:2 diues habebat oues et boues pluri- 12:2 The rich man had exceeding many flocks and herds: mos ualde 12:3 pauper autem nihil habebat omnino prae- 12:3 But the poor man had nothing, save one little ewe

lamb, which he had bought and nourished up: and it grew up together with him, and with his children; it did eat of his own meat, and drank of his own cup, and lay in his bosom, and was unto him as a daughter.

ter ouem unam paruulam quam emerat et nutrierat et quae creuerat apud eum cum filiis eius simul de pane illius comedens et de calice eius bibens et in sinu illius dormiens

Biblia Sacra Vulgata

12:4 And there came a traveller unto the rich man, and he eratque illi sicut filia 12:4 cum autem peregrinus quidam

spared to take of his own flock and of his own herd, to dress for the wayfaring man that was come unto him; but took the poor man’s lamb, and dressed it for the man that was come to him.

uenisset ad diuitem parcens ille sumere de ouibus et de bubus suis ut exhiberet conuiuium peregrino illi qui uenerat ad se tulit ouem uiri pauperis et praeparauit cibos homini

12:5 And David’s anger was greatly kindled against the qui uenerat ad se 12:5 iratus autem indignatione Dauid ad-

man; and he said to Nathan, As the LORD liveth, the man that hath done this thing shall surely die:

uersus hominem illum nimis dixit ad Nathan uiuit Domi-

12:6 And he shall restore the lamb fourfold, because he nus quoniam filius mortis est uir qui fecit hoc 12:6 ouem did this thing, and because he had no pity.

reddet in quadruplum eo quod fecerit uerbum istud et non

12:7 And Nathan said to David, Thou art the man. Thus pepercerit 12:7 dixit autem Nathan ad Dauid tu es ille uir

saith the LORD God of Israel, I anointed thee king over Israel, and I delivered thee out of the hand of Saul;

haec dicit Dominus Deus Israhel ego unxi te in regem su-

12:8 And I gave thee thy master’s house, and thy master’s per Israhel et ego erui te de manu Saul 12:8 et dedi tibi wives into thy bosom, and gave thee the house of Israel and of Judah; and if that had been too little, I would more-

over have given unto thee such and such things. domum domini tui et uxores domini tui in sinu tuo dedi-

que tibi domum Israhel et Iuda et si parua sunt ista adiciam

12:9 Wherefore hast thou despised the commandment of tibi multo maiora 12:9 quare ergo contempsisti uerbum Do-

the LORD, to do evil in his sight? thou hast killed Uriah the Hittite with the sword, and hast taken his wife to be thy wife, and hast slain him with the sword of the children of Ammon.

mini ut faceres malum in conspectu meo Vriam Hettheum percussisti gladio et uxorem illius accepisti uxorem et in-

12:10 Now therefore the sword shall never depart from terfecisti eum gladio filiorum Ammon 12:10 quam ob rem thine house; because thou hast despised me, and hast taken the wife of Uriah the Hittite to be thy wife.

non recedet gladius de domo tua usque in sempiternum eo quod despexeris me et tuleris uxorem Vriae Hetthei ut esset

12:11 Thus saith the LORD, Behold, I will raise up evil uxor tua 12:11 itaque haec dicit Dominus ecce ego susci-

against thee out of thine own house, and I will take thy wives before thine eyes, and give them unto thy neighbour, and he shall lie with thy wives in the sight of this sun. tabo super te malum de domo tua et tollam uxores tuas in

oculis tuis et dabo proximo tuo et dormiet cum uxoribus tuis

12:12 For thou didst it secretly: but I will do this thing in oculis solis huius 12:12 tu enim fecisti abscondite ego

before all Israel, and before the sun.

uero faciam uerbum istud in conspectu omnis Israhel et in

12:13 And David said unto Nathan, I have sinned against conspectu solis 12:13 et dixit Dauid ad Nathan peccaui Do-

the LORD. And Nathan said unto David, The LORD also hath put away thy sin; thou shalt not die.

mino dixitque Nathan ad Dauid Dominus quoque transtulit

12:14 Howbeit, because by this deed thou hast given great peccatum tuum non morieris 12:14 uerumtamen quoniam occasion to the enemies of the LORD to blaspheme, the child also that is born unto thee shall surely die.

blasphemare fecisti inimicos Domini propter uerbum hoc fi-

12:15 And Nathan departed unto his house. And the lius qui natus est tibi morte morietur 12:15 et reuersus est

LORD struck the child that Uriah’s wife bare unto David, and it was very sick.

Nathan domum suam percussitque Dominus paruulum quem

12:16 David therefore besought God for the child; and pepererat uxor Vriae Dauid et desperatus est 12:16 depre-

David fasted, and went in, and lay all night upon the earth.

catusque est Dauid Dominum pro paruulo et ieiunauit Da-

12:17 And the elders of his house arose, and went to him, uid ieiunio et ingressus seorsum iacuit super terram 12:17

to raise him up from the earth: but he would not, neither did he eat bread with them.

Liber II Samuelis

uenerunt autem seniores domus eius cogentes eum ut surge-

ret de terra qui noluit neque comedit cum eis cibum 12:18 12:18 And it came to pass on the seventh day, that the child died. And the servants of David feared to tell him that the child was dead: for they said, Behold, while the child was yet alive, we spake unto him, and he would not hearken unto our voice: how will he then vex himself, if we tell him that the child is dead?

accidit autem die septima ut moreretur infans timueruntque serui Dauid nuntiare ei quod mortuus esset paruulus dixerunt enim ecce cum paruulus adhuc uiueret loquebamur ad eum et non audiebat uocem nostram quanto magis si dixerimus mortuus est puer se adfliget 12:19 cum ergo uidisset Dauid 12:19 But when David saw that his servants whispered,

David perceived that the child was dead: therefore David said unto his servants, Is the child dead? And they said,

seruos suos musitantes intellexit quod mortuus esset infantu- He is dead. lus dixitque ad seruos suos num mortuus est puer qui responderunt ei mortuus est 12:20 surrexit igitur Dauid de terra et 12:20 Then David arose from the earth, and washed, and

anointed himself, and changed his apparel, and came into the house of the LORD, and worshipped: then he came to his own house; and when he required, they set bread before him, and he did eat.

lotus unctusque est cumque mutasset uestem ingressus est domum Domini et adorauit et uenit in domum suam petiuitque ut ponerent ei panem et comedit 12:21 dixerunt autem 12:21 Then said his servants unto him, What thing is this

that thou hast done? thou didst fast and weep for the child, while it was alive; but when the child was dead, thou didst

ei serui sui quis est sermo quem fecisti propter infantem cum rise and eat bread. adhuc uiueret ieiunasti et flebas mortuo autem puero surrexisti et comedisti panem 12:22 qui ait propter infantem 12:22 And he said, While the child was yet alive, I fasted

and wept: for I said, Who can tell whether GOD will be gracious to me, that the child may live?

dum adhuc uiueret ieiunaui et fleui dicebam enim quis scit si forte donet eum mihi Dominus et uiuet infans 12:23 nunc 12:23 But now he is dead, wherefore should I fast? can I

bring him back again? I shall go to him, but he shall not return to me.

autem quia mortuus est quare ieiuno numquid potero reuocare eum amplius ego uadam magis ad eum ille uero non reuertetur ad me 12:24 et consolatus est Dauid Bethsabee 12:24 And David comforted Bathsheba his wife, and went

in unto her, and lay with her: and she bare a son, and he called his name Solomon: and the LORD loved him.

uxorem suam ingressusque ad eam dormiuit cum ea quae genuit filium et uocauit nomen eius Salomon et Dominus dilexit eum 12:25 misitque in manu Nathan prophetae et uo- 12:25 And he sent by the hand of Nathan the prophet; and

he called his name Jedidiah, because of the LORD.

cauit nomen eius Amabilis Domino eo quod diligeret eum Dominus 12:26 igitur pugnabat Ioab contra Rabbath filio- 12:26 And Joab fought against Rabbah of the children of

Ammon, and took the royal city.

rum Ammon et expugnabat urbem regiam 12:27 misitque 12:27 And Joab sent messengers to David, and said, I have

fought against Rabbah, and have taken the city of waters.

Ioab nuntios ad Dauid dicens dimicaui aduersum Rabbath

et capienda est urbs Aquarum 12:28 nunc igitur congrega 12:28 Now therefore gather the rest of the people together, and encamp against the city, and take it: lest I take the city, and it be called after my name.

reliquam partem populi et obside ciuitatem et cape eam ne cum a me uastata fuerit urbs nomini meo adscribatur uictoria 12:29 congregauit itaque Dauid omnem populum et pro- 12:29 And David gathered all the people together, and

went to Rabbah, and fought against it, and took it.

fectus est aduersum Rabbath cumque dimicasset cepit eam

Biblia Sacra Vulgata

12:30 And he took their king’s crown from off his head, 12:30 et tulit diadema regis eorum de capite eius pondo auri

the weight whereof was a talent of gold with the precious stones: and it was set on David’s head. And he brought

forth the spoil of the city in great abundance. talentum habens gemmas pretiosissimas et inpositum est su-

per caput Dauid sed et praedam ciuitatis asportauit multam

12:31 And he brought forth the people that were therein, ualde 12:31 populum quoque eius adducens serrauit et cir-

and put them under saws, and under harrows of iron, and under axes of iron, and made them pass through the brickkiln: and thus did he unto all the cities of the children of Ammon. So David and all the people returned unto Jerusalem.

cumegit super eos ferrata carpenta diuisitque cultris et transduxit in typo laterum sic fecit uniuersis ciuitatibus filiorum Ammon et reuersus est Dauid et omnis exercitus Hierusalem

13:1 And it came to pass after this, that Absalom the son 13:1 factum est autem post haec ut Absalom filii Dauid so-

of David had a fair sister, whose name was Tamar; and Amnon the son of David loved her.

rorem speciosissimam uocabulo Thamar adamaret Amnon

13:2 And Amnon was so vexed, that he fell sick for his filius Dauid 13:2 et deperiret eam ualde ita ut aegrotaret

sister Tamar; for she was a virgin; and Amnon thought it hard for him to do anything to her.

propter amorem eius quia cum esset uirgo difficile ei uide-

13:3 But Amnon had a friend, whose name was Jonadab, batur ut quippiam inhoneste ageret cum ea 13:3 erat autem

the son of Shimeah David’s brother: and Jonadab was a very subtil man.

Amnonis amicus nomine Ionadab filius Semaa fratris Da-

13:4 And he said unto him, Why art thou, being the king’s uid uir prudens ualde 13:4 qui dixit ad eum quare sic adson, lean from day to day? wilt thou not tell me? And Amnon said unto him, I love Tamar, my brother Absalom’s

sister. tenuaris macie fili regis per singulos dies cur non indicas

mihi dixitque ei Amnon Thamar sororem Absalom fratris

13:5 And Jonadab said unto him, Lay thee down on thy mei amo 13:5 cui respondit Ionadab cuba super lectulum

bed, and make thyself sick: and when thy father cometh to see thee, say unto him, I pray thee, let my sister Tamar come, and give me meat, and dress the meat in my sight, that I may see it, and eat it at her hand.

tuum et languorem simula cumque uenerit pater tuus ut uisitet te dic ei ueniat oro Thamar soror mea ut det mihi cibum

13:6 So Amnon lay down, and made himself sick: and et faciat pulmentum ut comedam de manu eius 13:6 accu-

when the king was come to see him, Amnon said unto the king, I pray thee, let Tamar my sister come, and make me

a couple of cakes in my sight, that I may eat at her hand. buit itaque Amnon et quasi aegrotare coepit cumque uenisset rex ad uisitandum eum ait Amnon ad regem ueniat obsecro Thamar soror mea ut faciat in oculis meis duas sorbitiuncu-

13:7 Then David sent home to Tamar, saying, Go now to las et cibum capiam de manu eius 13:7 misit ergo Dauid ad

thy brother Amnon’s house, and dress him meat.

Thamar domum dicens ueni in domum Amnon fratris tui et

13:8 So Tamar went to her brother Amnon’s house; and fac ei pulmentum 13:8 uenitque Thamar in domum Amnon

he was laid down. And she took flour, and kneaded it, and made cakes in his sight, and did bake the cakes.

fratris sui ille autem iacebat quae tollens farinam commis-

13:9 And she took a pan, and poured them out before him; cuit et liquefaciens in oculis eius coxit sorbitiunculas 13:9

but he refused to eat. And Amnon said, Have out all men from me. And they went out every man from him.

tollensque quod coxerat effudit et posuit coram eo et noluit comedere dixitque Amnon eicite uniuersos a me cum-

13:10 And Amnon said unto Tamar, Bring the meat into que eiecissent omnes 13:10 dixit Amnon ad Thamar infer

the chamber, that I may eat of thine hand. And Tamar took the cakes which she had made, and brought them into the

chamber to Amnon her brother. cibum in conclaue ut uescar de manu tua tulit ergo Thamar sorbitiunculas quas fecerat et intulit ad Amnon fratrem suum

Liber II Samuelis

in conclaue 13:11 cumque obtulisset ei cibum adprehendit 13:11 And when she had brought them unto him to eat, he

took hold of her, and said unto her, Come lie with me, my sister.

eam et ait ueni cuba mecum soror mea 13:12 quae respon-

13:12 And she answered him, Nay, my brother, do not force me; for no such thing ought to be done in Israel: do not thou this folly.

dit ei noli frater mi noli opprimere me neque enim hoc fas est in Israhel noli facere stultitiam hanc 13:13 et ego enim 13:13 And I, whither shall I cause my shame to go? and

as for thee, thou shalt be as one of the fools in Israel. Now therefore, I pray thee, speak unto the king; for he will not

ferre non potero obprobrium meum et tu eris quasi unus withhold me from thee. de insipientibus in Israhel quin potius loquere ad regem et non negabit me tibi 13:14 noluit autem adquiescere preci- 13:14 Howbeit he would not hearken unto her voice: but,

being stronger than she, forced her, and lay with her.

bus eius sed praeualens uiribus oppressit eam et cubauit cum illa 13:15 et exosam eam habuit Amnon magno odio nimis 13:15 Then Amnon hated her exceedingly; so that the

hatred wherewith he hated her was greater than the love wherewith he had loved her. And Amnon said unto her,

ita ut maius esset odium quo oderat eam amore quo ante di- Arise, be gone. lexerat dixitque ei Amnon surge uade 13:16 quae respondit 13:16 And she said unto him, There is no cause: this evil

in sending me away is greater than the other that thou didst unto me. But he would not hearken unto her.

ei maius est hoc malum quod nunc agis aduersum me quam

quod ante fecisti expellens me et noluit audire eam 13:17 13:17 Then he called his servant that ministered unto him, and said, Put now this woman out from me, and bolt the door after her.

sed uocato puero qui ministrabat ei dixit eice hanc a me foras et claude ostium post eam 13:18 quae induta erat talari 13:18 And she had a garment of divers colours upon her:

for with such robes were the king’s daughters that were virgins apparelled. Then his servant brought her out, and

tunica huiuscemodi enim filiae regis uirgines uestibus ute- bolted the door after her. bantur eiecit itaque eam minister illius foras clausitque fores post eam 13:19 quae aspergens cinerem capiti suo scissa 13:19 And Tamar put ashes on her head, and rent her gar-

ment of divers colours that was on her, and laid her hand on her head, and went on crying.

talari tunica inpositisque manibus super caput suum ibat ingrediens et clamans 13:20 dixit autem ei Absalom frater 13:20 And Absalom her brother said unto her, Hath Am-

non thy brother been with thee? but hold now thy peace, my sister: he is thy brother; regard not this thing. So

suus num Amnon frater tuus concubuit tecum sed nunc so- Tamar remained desolate in her brother Absalom’s house.

ror tace frater tuus est neque adfligas cor tuum pro re hac mansit itaque Thamar contabescens in domo Absalom fratris sui 13:21 cum autem audisset rex Dauid uerba haec con- 13:21 But when king David heard of all these things, he

was very wroth.

tristatus est ualde 13:22 porro non est locutus Absalom ad 13:22 And Absalom spake unto his brother Amnon neither

good nor bad: for Absalom hated Amnon, because he had forced his sister Tamar.

Amnon nec malum nec bonum oderat enim Absalom Amnon eo quod uiolasset Thamar sororem suam 13:23 factum 13:23 And it came to pass after two full years, that Ab-

salom had sheepshearers in Baalhazor, which is beside Ephraim: and Absalom invited all the king’s sons.

est autem post tempus biennii ut tonderentur oues Absalom in Baalasor quae est iuxta Ephraim et uocauit Absalom omnes filios regis 13:24 uenitque ad regem et ait ad eum ecce 13:24 And Absalom came to the king, and said, Behold

now, thy servant hath sheepshearers; let the king, I beseech thee, and his servants go with thy servant.

tondentur oues serui tui ueniat oro rex cum seruis suis ad seruum suum 13:25 dixitque rex ad Absalom noli fili mi 13:25 And the king said to Absalom, Nay, my son, let us

not all now go, lest we be chargeable unto thee. And he pressed him: howbeit he would not go, but blessed him.

noli rogare ut ueniamus omnes et grauemus te cum autem

Biblia Sacra Vulgata

13:26 Then said Absalom, If not, I pray thee, let my cogeret eum et noluisset ire benedixit ei 13:26 et ait Absa-

brother Amnon go with us. And the king said unto him, Why should he go with thee?

lom si non uis uenire ueniat obsecro nobiscum saltem Amnon frater meus dixitque ad eum rex non est necesse ut ua-

13:27 But Absalom pressed him, that he let Amnon and dat tecum 13:27 coegit itaque eum Absalom et dimisit cum

all the king’s sons go with him.

13:28 Now Absalom had commanded his servants, saying, eo Amnon et uniuersos filios regis 13:28 praeceperat autem

Mark ye now when Amnon’s heart is merry with wine, and when I say unto you, Smite Amnon; then kill him, fear not: have not I commanded you? be courageous, and be valiant.

Absalom pueris suis dicens obseruate cum temulentus fuerit Amnon uino et dixero uobis percutite eum et interficite nolite timere ego enim sum qui praecepi uobis roboramini

13:29 And the servants of Absalom did unto Amnon as et estote uiri fortes 13:29 fecerunt ergo pueri Absalom ad-

Absalom had commanded. Then all the king’s sons arose, and every man gat him up upon his mule, and fled.

uersum Amnon sicut praeceperat eis Absalom surgentesque omnes filii regis ascenderunt singuli mulas suas et fugerunt

13:30 And it came to pass, while they were in the way, that 13:30 cumque adhuc pergerent in itinere fama praeuenit ad

tidings came to David, saying, Absalom hath slain all the king’s sons, and there is not one of them left.

Dauid dicens percussit Absalom omnes filios regis et non re-

13:31 Then the king arose, and tare his garments, and lay mansit ex eis saltem unus 13:31 surrexit itaque rex et scidit

on the earth; and all his servants stood by with their clothes rent.

uestimenta sua et cecidit super terram et omnes serui ipsius

13:32 And Jonadab, the son of Shimeah David’s brother, qui adsistebant ei sciderunt uestimenta sua 13:32 respon-

answered and said, Let not my lord suppose that they have slain all the young men the king’s sons; for Amnon only is dead: for by the appointment of Absalom this hath been determined from the day that he forced his sister Tamar.

dens autem Ionadab filius Samaa fratris Dauid dixit ne aestimet dominus meus quod omnes pueri filii regis occisi sint Amnon solus mortuus est quoniam in ore Absalom erat po-

13:33 Now therefore let not my lord the king take the thing situs ex die qua oppressit Thamar sororem eius 13:33 nunc to his heart, to think that all the king’s sons are dead: for Amnon only is dead.

ergo ne ponat dominus meus rex super cor suum uerbum istud dicens omnes filii regis occisi sunt quoniam Amnon so-

13:34 But Absalom fled. And the young man that kept the lus mortuus est 13:34 fugit autem Absalom et leuauit puer

watch lifted up his eyes, and looked, and, behold, there came much people by the way of the hill side behind him.

speculator oculos suos et aspexit et ecce populus multus ue-

13:35 And Jonadab said unto the king, Behold, the king’s niebat per iter deuium ex latere montis 13:35 dixit autem

sons come: as thy servant said, so it is.

Ionadab ad regem ecce filii regis adsunt iuxta uerbum serui

13:36 And it came to pass, as soon as he had made an tui sic factum est 13:36 cumque cessasset loqui apparuerunt

end of speaking, that, behold, the king’s sons came, and lifted up their voice and wept: and the king also and all

his servants wept very sore. et filii regis et intrantes leuauerunt uocem suam et fleuerunt

sed et rex et omnes serui eius fleuerunt ploratu magno nimis

13:37 But Absalom fled, and went to Talmai, the son of 13:37 porro Absalom fugiens abiit ad Tholomai filium Amiur

Ammihud, king of Geshur. And David mourned for his son every day.

regem Gessur luxit ergo Dauid filium suum cunctis diebus

13:38 So Absalom fled, and went to Geshur, and was there 13:38 Absalom autem cum fugisset et uenisset in Gessur fuit

three years.

13:39 And the soul of king David longed to go forth unto ibi tribus annis 13:39 cessauitque Dauid rex persequi Absa-

Absalom: for he was comforted concerning Amnon, seeing he was dead.

Liber II Samuelis

lom eo quod consolatus esset super Amnon interitu

14:1 intellegens autem Ioab filius Saruiae quod cor regis 14:1 Now Joab the son of Zeruiah perceived that the king’s

heart was toward Absalom.

uersum esset ad Absalom 14:2 misit Thecuam et tulit inde 14:2 And Joab sent to Tekoah, and fetched thence a wise

woman, and said unto her, I pray thee, feign thyself to be a mourner, and put on now mourning apparel, and anoint not thyself with oil, but be as a woman that had a long time mourned for the dead:

mulierem sapientem dixitque ad eam lugere te simula et induere ueste lugubri et ne unguaris oleo ut sis quasi mulier plurimo iam tempore lugens mortuum 14:3 et ingredieris 14:3 And come to the king, and speak on this manner unto

him. So Joab put the words in her mouth.

ad regem et loqueris ad eum sermones huiuscemodi posuit autem Ioab uerba in ore eius 14:4 itaque cum ingressa fuis- 14:4 And when the woman of Tekoah spake to the king,

she fell on her face to the ground, and did obeisance, and said, Help, O king.

set mulier thecuites ad regem cecidit coram eo super terram et adorauit et dixit serua me rex 14:5 et ait ad eam rex 14:5 And the king said unto her, What aileth thee? And

she answered, I am indeed a widow woman, and mine husband is dead.

quid causae habes quae respondit heu mulier uidua ego sum mortuus est enim uir meus 14:6 et ancillae tuae erant duo 14:6 And thy handmaid had two sons, and they two strove

together in the field, and there was none to part them, but the one smote the other, and slew him.

filii qui rixati sunt aduersum se in agro nullusque erat qui eos prohibere posset et percussit alter alterum et interfecit

eum 14:7 et ecce consurgens uniuersa cognatio aduersum 14:7 And, behold, the whole family is risen against thine handmaid, and they said, Deliver him that smote his brother, that we may kill him, for the life of his brother whom he slew; and we will destroy the heir also: and so they shall quench my coal which is left, and shall not leave to my husband neither name nor remainder upon the earth.

ancillam tuam dicit trade eum qui percussit fratrem suum ut occidamus eum pro anima fratris sui quem interfecit et deleamus heredem et quaerunt extinguere scintillam meam quae relicta est ut non supersit uiro meo nomen et reliquiae super

terram 14:8 et ait rex ad mulierem uade in domum tuam 14:8 And the king said unto the woman, Go to thine house,

and I will give charge concerning thee.

et ego iubebo pro te 14:9 dixitque mulier thecuites ad re- 14:9 And the woman of Tekoah said unto the king, My

lord, O king, the iniquity be on me, and on my father’s house: and the king and his throne be guiltless.

gem in me domine mi rex iniquitas et in domum patris mei rex autem et thronus eius sit innocens 14:10 et ait rex qui 14:10 And the king said, Whoever saith ought unto thee,

bring him to me, and he shall not touch thee any more.

contradixerit tibi adduc eum ad me et ultra non addet ut tangat te 14:11 quae ait recordetur rex Domini Dei sui ut non 14:11 Then said she, I pray thee, let the king remember

the LORD thy God, that thou wouldest not suffer the revengers of blood to destroy any more, lest they destroy my son. And he said, As the LORD liveth, there shall not one hair of thy son fall to the earth.

multiplicentur proximi sanguinis ad ulciscendum et nequaquam interficient filium meum qui ait uiuit Dominus quia non cadet de capillis filii tui super terram 14:12 dixit ergo 14:12 Then the woman said, Let thine handmaid, I pray

thee, speak one word unto my lord the king. And he said, Say on.

mulier loquatur ancilla tua ad dominum meum regem uerbum et ait loquere 14:13 dixitque mulier quare cogitasti is- 14:13 And the woman said, Wherefore then hast thou

thought such a thing against the people of God? for the king doth speak this thing as one which is faulty, in that

tiusmodi rem contra populum Dei et locutus est rex uerbum the king doth not fetch home again his banished. istud ut peccet et non reducat eiectum suum 14:14 omnes 14:14 For we must needs die, and are as water spilt on the

ground, which cannot be gathered up again; neither doth God respect any person: yet doth he devise means, that his

morimur et quasi aquae delabimur in terram quae non re- banished be not expelled from him.

Biblia Sacra Vulgata

uertuntur nec uult perire Deus animam sed retractat cogitans

14:15 Now therefore that I am come to speak of this thing ne penitus pereat qui abiectus est 14:15 nunc igitur ueni ut

unto my lord the king, it is because the people have made me afraid: and thy handmaid said, I will now speak unto the king; it may be that the king will perform the request of his handmaid.

loquar ad regem dominum meum uerbum hoc praesente populo et dixit ancilla tua loquar ad regem si quo modo faciat

14:16 For the king will hear, to deliver his handmaid out rex uerbum ancillae suae 14:16 et audiuit rex ut liberaret

of the hand of the man that would destroy me and my son together out of the inheritance of God.

ancillam suam de manu omnium qui uolebant delere me et

14:17 Then thine handmaid said, The word of my lord the filium meum simul de hereditate Dei 14:17 dicat ergo an-

king shall now be comfortable: for as an angel of God, so is my lord the king to discern good and bad: therefore the

LORD thy God will be with thee. cilla tua ut fiat uerbum domini mei regis quasi sacrificium sicut enim angelus Dei sic est dominus meus rex ut nec benedictione nec maledictione moueatur unde et Dominus De-

14:18 Then the king answered and said unto the woman, us tuus est tecum 14:18 et respondens rex dixit ad mulie-

Hide not from me, I pray thee, the thing that I shall ask thee. And the woman said, Let my lord the king now

speak. rem ne abscondas a me uerbum quod te interrogo dixitque 14:19 And the king said, Is not the hand of Joab with thee mulier loquere domine mi rex 14:19 et ait rex numquid ma-

in all this? And the woman answered and said, As thy soul liveth, my lord the king, none can turn to the right hand or to the left from ought that my lord the king hath spoken: for thy servant Joab, he bade me, and he put all these words in the mouth of thine handmaid:

nus Ioab tecum est in omnibus istis respondit mulier et ait per salutem animae tuae domine mi rex nec ad dextram nec ad sinistram est ex omnibus his quae locutus est dominus meus rex seruus enim tuus Ioab ipse praecepit mihi et ipse

14:20 To fetch about this form of speech hath thy servant posuit in os ancillae tuae omnia uerba haec 14:20 ut uerte-

Joab done this thing: and my lord is wise, according to the wisdom of an angel of God, to know all things that are in

the earth. rem figuram sermonis huius seruus tuus Ioab praecepit istud

tu autem domine mi sapiens es sicut habet sapientiam an-

14:21 And the king said unto Joab, Behold now, I have gelus Dei ut intellegas omnia super terram 14:21 et ait rex

done this thing: go therefore, bring the young man Absalom again.

ad Ioab ecce placatus feci uerbum tuum uade igitur et re-

14:22 And Joab fell to the ground on his face, and bowed uoca puerum Absalom 14:22 cadensque Ioab super faciem

himself, and thanked the king: and Joab said, To day thy servant knoweth that I have found grace in thy sight, my lord, O king, in that the king hath fulfilled the request of his servant.

suam in terram adorauit et benedixit regi et dixit Ioab hodie intellexit seruus tuus quia inueni gratiam in oculis tuis do-

14:23 So Joab arose and went to Geshur, and brought Ab- mine mi rex fecisti enim sermonem serui tui 14:23 surrexit

salom to Jerusalem.

ergo Ioab et abiit in Gessur et adduxit Absalom in Hieru-

14:24 And the king said, Let him turn to his own house, salem 14:24 dixit autem rex reuertatur in domum suam et

and let him not see my face. So Absalom returned to his own house, and saw not the king’s face.

faciem meam non uideat reuersus est itaque Absalom in do-

14:25 But in all Israel there was none to be so much mum suam et faciem regis non uidit 14:25 porro sicut Ab-

praised as Absalom for his beauty: from the sole of his foot even to the crown of his head there was no blemish in

him. salom uir non erat pulcher in omni Israhel et decorus nimis

a uestigio pedis usque ad uerticem non erat in eo ulla ma-

14:26 And when he polled his head, (for it was at every cula 14:26 et quando tondebatur capillum semel autem in

year’s end that he polled it: because the hair was heavy on him, therefore he polled it:) he weighed the hair of his head at two hundred shekels after the king’s weight.

Liber II Samuelis

anno tondebatur quia grauabat eum caesaries ponderabat capillos capitis sui ducentis siclis pondere publico 14:27 nati 14:27 And unto Absalom there were born three sons, and

one daughter, whose name was Tamar: she was a woman of a fair countenance.

sunt autem Absalom filii tres et filia una nomine Thamar eleganti forma 14:28 mansitque Absalom Hierusalem duobus 14:28 So Absalom dwelt two full years in Jerusalem, and

saw not the king’s face.

annis et faciem regis non uidit 14:29 misit itaque ad Ioab 14:29 Therefore Absalom sent for Joab, to have sent him

to the king; but he would not come to him: and when he sent again the second time, he would not come.

ut mitteret eum ad regem qui noluit uenire ad eum cumque secundo misisset et ille noluisset uenire 14:30 dixit seruis 14:30 Therefore he said unto his servants, See, Joab’s field

is near mine, and he hath barley there; go and set it on fire. And Absalom’s servants set the field on fire.

suis scitis agrum Ioab iuxta agrum meum habentem messem hordei ite igitur et succendite eum igni succenderunt ergo serui Absalom segetem igni 14:31 surrexitque Ioab et ue- 14:31 Then Joab arose, and came to Absalom unto his

house, and said unto him, Wherefore have thy servants set my field on fire?

nit ad Absalom in domum eius et dixit quare succenderunt serui tui segetem meam igni 14:32 et respondit Absalom ad 14:32 And Absalom answered Joab, Behold, I sent unto

thee, saying, Come hither, that I may send thee to the king, to say, Wherefore am I come from Geshur? it had been good for me to have been there still: now therefore let me see the king’s face; and if there be any iniquity in me, let him kill me.

Ioab misi ad te obsecrans ut uenires ad me et mitterem te ad regem ut diceres ei quare ueni de Gessur melius mihi erat ibi esse obsecro ergo ut uideam faciem regis quod si memor est iniquitatis meae interficiat me 14:33 ingressus Ioab ad 14:33 So Joab came to the king, and told him: and when he

had called for Absalom, he came to the king, and bowed himself on his face to the ground before the king: and the

regem nuntiauit ei uocatusque Absalom intrauit ad regem et king kissed Absalom. adorauit super faciem terrae coram eo osculatusque est rex Absalom

15:1 igitur post haec fecit sibi Absalom currum et equites 15:1 And it came to pass after this, that Absalom prepared

him chariots and horses, and fifty men to run before him.

et quinquaginta uiros qui praecederent eum 15:2 et mane 15:2 And Absalom rose up early, and stood beside the way

of the gate: and it was so, that when any man that had a controversy came to the king for judgment, then Absalom called unto him, and said, Of what city art thou? And he said, Thy servant is of one of the tribes of Israel.

consurgens Absalom stabat iuxta introitum portae et omnem uirum qui habebat negotium ut ueniret ad regis iudicium uocabat Absalom ad se et dicebat de qua ciuitate es tu qui respondens aiebat ex una tribu Israhel ego sum seruus tuus 15:3 respondebatque ei Absalom uidentur mihi sermones tui 15:3 And Absalom said unto him, See, thy matters are

good and right; but there is no man deputed of the king to hear thee.

boni et iusti sed non est qui te audiat constitutus a rege dicebatque Absalom 15:4 quis me constituat iudicem super 15:4 Absalom said moreover, Oh that I were made judge

in the land, that every man which hath any suit or cause might come unto me, and I would do him justice!

terram ut ad me ueniant omnes qui habent negotium et iuste iudicem 15:5 sed et cum accederet ad eum homo ut salu- 15:5 And it was so, that when any man came nigh to him

to do him obeisance, he put forth his hand, and took him, and kissed him.

taret illum extendebat manum suam et adprehendens osculabatur eum 15:6 faciebatque hoc omni Israhel qui uenie- 15:6 And on this manner did Absalom to all Israel that

came to the king for judgment: so Absalom stole the hearts of the men of Israel.

bat ad iudicium ut audiretur a rege et sollicitabat corda uiro-

Biblia Sacra Vulgata

15:7 And it came to pass after forty years, that Absalom rum Israhel 15:7 post quattuor autem annos dixit Absalom

said unto the king, I pray thee, let me go and pay my vow, which I have vowed unto the LORD, in Hebron.

ad regem uadam et reddam uota mea quae uoui Domino in

15:8 For thy servant vowed a vow while I abode at Geshur Hebron 15:8 uouens enim uouit seruus tuus cum esset in in Syria, saying, If the LORD shall bring me again indeed to Jerusalem, then I will serve the LORD.

Gessur Syriae dicens si reduxerit me Dominus in Hierusa-

15:9 And the king said unto him, Go in peace. So he arose, lem sacrificabo Domino 15:9 dixitque ei rex uade in pace et

and went to Hebron. 15:10 But Absalom sent spies throughout all the tribes of surrexit et abiit in Hebron 15:10 misit autem Absalom ex-Israel, saying, As soon as ye hear the sound of the trumpet, then ye shall say, Absalom reigneth in Hebron.

ploratores in uniuersas tribus Israhel dicens statim ut audieritis clangorem bucinae dicite regnauit Absalom in Hebron

15:11 And with Absalom went two hundred men out of 15:11 porro cum Absalom ierunt ducenti uiri de Hierusalem

Jerusalem, that were called; and they went in their simplicity, and they knew not any thing.

uocati euntes simplici corde et causam penitus ignorantes

15:12 And Absalom sent for Ahithophel the Gilonite, 15:12 accersiuit quoque Absalom Ahitofel Gilonitem consi-

David’s counsellor, from his city, even from Giloh, while he offered sacrifices. And the conspiracy was strong; for

the people increased continually with Absalom. liarium Dauid de ciuitate sua Gilo cum immolaret uictimas

et facta est coniuratio ualida populusque concurrens augeba-

15:13 And there came a messenger to David, saying, The tur cum Absalom 15:13 uenit igitur nuntius ad Dauid dicens

hearts of the men of Israel are after Absalom.

15:14 And David said unto all his servants that were with toto corde uniuersus Israhel sequitur Absalom 15:14 et ait

him at Jerusalem, Arise, and let us flee; for we shall not else escape from Absalom: make speed to depart, lest he overtake us suddenly, and bring evil upon us, and smite the city with the edge of the sword.

Dauid seruis suis qui erant cum eo in Hierusalem surgite fugiamus neque enim erit nobis effugium a facie Absalom festinate egredi ne forte ueniens occupet nos et inpellat super

15:15 And the king’s servants said unto the king, Behold, nos ruinam et percutiat ciuitatem in ore gladii 15:15 dixe-

thy servants are ready to do whatsoever my lord the king shall appoint.

runtque serui regis ad eum omnia quaecumque praeceperit

15:16 And the king went forth, and all his household after dominus noster rex libenter exsequimur serui tui 15:16 eg-

him. And the king left ten women, which were concubines, to keep the house.

ressus est ergo rex et uniuersa domus eius pedibus suis et dereliquit rex decem mulieres concubinas ad custodiendam

15:17 And the king went forth, and all the people after domum 15:17 egressusque rex et omnis Israhel pedibus suis

him, and tarried in a place that was far off.

15:18 And all his servants passed on beside him; and all stetit procul a domo 15:18 et uniuersi serui eius ambulabant

the Cherethites, and all the Pelethites, and all the Gittites, six hundred men which came after him from Gath, passed

on before the king. iuxta eum et legiones Cherethi et Felethi et omnes Getthei

sescenti uiri qui secuti eum fuerant de Geth praecedebant

15:19 Then said the king to Ittai the Gittite, Wherefore regem 15:19 dixit autem rex ad Ethai Gettheum cur uenis goest thou also with us? return to thy place, and abide with the king: for thou art a stranger, and also an exile.

nobiscum reuertere et habita cum rege quia peregrinus es et

15:20 Whereas thou camest but yesterday, should I this egressus de loco tuo 15:20 heri uenisti et hodie inpelleris

day make thee go up and down with us? seeing I go whither I may, return thou, and take back thy brethren:

mercy and truth be with thee. nobiscum egredi ego autem uadam quo iturus sum reuertere 15:21 And Ittai answered the king, and said, As the LORD et reduc tecum fratres tuos ostendisti gratiam et fidem 15:21

liveth, and as my lord the king liveth, surely in what place my lord the king shall be, whether in death or life, even

there also will thy servant be. et respondit Ethai regi dicens uiuit Dominus et uiuit domi-

Liber II Samuelis

nus meus rex quoniam in quocumque loco fueris domine mi rex siue in morte siue in uita ibi erit seruus tuus 15:22 et ait 15:22 And David said to Ittai, Go and pass over. And Ittai

the Gittite passed over, and all his men, and all the little ones that were with him.

Dauid Ethai ueni et transi et transiuit Ethai Gettheus et omnes uiri qui cum eo erant et reliqua multitudo 15:23 omnes- 15:23 And all the country wept with a loud voice, and all

the people passed over: the king also himself passed over the brook Kidron, and all the people passed over, toward

que flebant uoce magna et uniuersus populus transiebat rex the way of the wilderness. quoque transgrediebatur torrentem Cedron et cunctus populus incedebat contra uiam quae respicit ad desertum 15:24 15:24 And lo Zadok also, and all the Levites were with

him, bearing the ark of the covenant of God: and they set down the ark of God; and Abiathar went up, until all the

uenit autem et Sadoc et uniuersi Leuitae cum eo portantes ar- people had done passing out of the city. cam foederis Dei et deposuerunt arcam Dei et ascendit Abiathar donec expletus est omnis populus qui egressus fuerat de ciuitate 15:25 et dixit rex ad Sadoc reporta arcam Dei 15:25 And the king said unto Zadok, Carry back the ark of

God into the city: if I shall find favour in the eyes of the LORD, he will bring me again, and shew me both it, and

in urbem si inuenero gratiam in oculis Domini reducet me his habitation:

et ostendet mihi eam et tabernaculum suum 15:26 si autem 15:26 But if he thus say, I have no delight in thee; behold,

here am I, let him do to me as seemeth good unto him.

dixerit non places praesto sum faciat quod bonum est coram se 15:27 et dixit rex ad Sadoc sacerdotem o uidens reuertere 15:27 The king said also unto Zadok the priest, Art not

thou a seer? return into the city in peace, and your two sons with you, Ahimaaz thy son, and Jonathan the son of

in ciuitatem in pace et Achimaas filius tuus et Ionathan filius Abiathar.

Abiathar duo filii uestri sint uobiscum 15:28 ecce ego ab- 15:28 See, I will tarry in the plain of the wilderness, until

there come word from you to certify me.

scondar in campestribus deserti donec ueniat sermo a uobis indicans mihi 15:29 reportauerunt igitur Sadoc et Abiathar 15:29 Zadok therefore and Abiathar carried the ark of God

again to Jerusalem: and they tarried there.

arcam Dei Hierusalem et manserunt ibi 15:30 porro Dauid 15:30 And David went up by the ascent of mount Olivet,

and wept as he went up, and had his head covered, and he went barefoot: and all the people that was with him covered every man his head, and they went up, weeping as they went up.

ascendebat cliuum Oliuarum scandens et flens operto capite et nudis pedibus incedens sed et omnis populus qui erat cum eo operto capite ascendebat plorans 15:31 nuntiatum est au- 15:31 And one told David, saying, Ahithophel is among

the conspirators with Absalom. And David said, O LORD, I pray thee, turn the counsel of Ahithophel into foolish-

tem Dauid quod et Ahitofel esset in coniuratione cum Absa- ness. lom dixitque Dauid infatua quaeso consilium Ahitofel Domine 15:32 cumque ascenderet Dauid summitatem montis 15:32 And it came to pass, that when David was come to

the top of the mount, where he worshipped God, behold, Hushai the Archite came to meet him with his coat rent,

in quo adoraturus erat Dominum ecce occurrit ei Husai Ara- and earth upon his head: chites scissa ueste et terra pleno capite 15:33 et dixit ei Da- 15:33 Unto whom David said, If thou passest on with me,

then thou shalt be a burden unto me:

uid si ueneris mecum eris mihi oneri 15:34 si autem in ciui- 15:34 But if thou return to the city, and say unto Absalom,

I will be thy servant, O king; as I have been thy father’s servant hitherto, so will I now also be thy servant: then

tatem reuertaris et dixeris Absalom seruus tuus sum rex sicut mayest thou for me defeat the counsel of Ahithophel. fui seruus patris tui sic ero seruus tuus dissipabis consilium Ahitofel 15:35 habes autem tecum Sadoc et Abiathar sacer- 15:35 And hast thou not there with thee Zadok and Abi-

athar the priests? therefore it shall be, that what thing soever thou shalt hear out of the king’s house, thou shalt tell

dotes et omne uerbum quodcumque audieris de domo regis it to Zadok and Abiathar the priests.

Biblia Sacra Vulgata

15:36 Behold, they have there with them their two sons, indicabis Sadoc et Abiathar sacerdotibus 15:36 sunt autem

Ahimaaz Zadok’s son, and Jonathan Abiathar’s son; and by them ye shall send unto me every thing that ye can

hear. cum eis duo filii eorum Achimaas Sadoc et Ionathan Abia-

thar et mittetis per eos ad me omne uerbum quod audieritis

15:37 So Hushai David’s friend came into the city, and 15:37 ueniente ergo Husai amico Dauid in ciuitatem Absalom

Absalom came into Jerusalem.

quoque ingressus est Hierusalem

16:1 And when David was a little past the top of the hill, 16:1 cumque Dauid transisset paululum montis uerticem

behold, Ziba the servant of Mephibosheth met him, with a couple of asses saddled, and upon them two hundred loaves of bread, and an hundred bunches of raisins, and an hundred of summer fruits, and a bottle of wine.

apparuit Siba puer Mifiboseth in occursum eius cum duobus asinis qui onerati erant ducentis panibus et centum alligaturis uuae passae et centum massis palatarum et utribus

16:2 And the king said unto Ziba, What meanest thou by uini 16:2 et dixit rex Sibae quid sibi uolunt haec respondit-

these? And Ziba said, The asses be for the king’s household to ride on; and the bread and summer fruit for the young men to eat; and the wine, that such as be faint in the wilderness may drink.

que Siba asini domestici regis ut sedeant et panes et palatae ad uescendum pueris tuis uinum autem ut bibat si quis

16:3 And the king said, And where is thy master’s son? defecerit in deserto 16:3 et ait rex ubi est filius domini tui

And Ziba said unto the king, Behold, he abideth at Jerusalem: for he said, To day shall the house of Israel

restore me the kingdom of my father. responditque Siba regi remansit in Hierusalem dicens hodie 16:4 Then said the king to Ziba, Behold, thine are all that restituet mihi domus Israhel regnum patris mei 16:4 et ait

pertained unto Mephibosheth. And Ziba said, I humbly beseech thee that I may find grace in thy sight, my lord, O

king. rex Sibae tua sint omnia quae fuerunt Mifiboseth dixitque 16:5 And when king David came to Bahurim, behold, Siba adoro inueniam gratiam coram te domine mi rex 16:5

thence came out a man of the family of the house of Saul, whose name was Shimei, the son of Gera: he came forth,

and cursed still as he came. uenit ergo rex Dauid usque Baurim et ecce egrediebatur inde

uir de cognatione domus Saul nomine Semei filius Gera pro-

16:6 And he cast stones at David, and at all the servants cedebat egrediens et maledicebat 16:6 mittebatque lapides

of king David: and all the people and all the mighty men were on his right hand and on his left.

contra Dauid et contra uniuersos seruos regis Dauid omnis autem populus et uniuersi bellatores a dextro et sinistro la-

16:7 And thus said Shimei when he cursed, Come out, tere regis incedebant 16:7 ita autem loquebatur Semei cum

come out, thou bloody man, and thou man of Belial:

malediceret regi egredere egredere uir sanguinum et uir Be-

16:8 The LORD hath returned upon thee all the blood of lial 16:8 reddidit tibi Dominus uniuersum sanguinem do-

the house of Saul, in whose stead thou hast reigned; and the LORD hath delivered the kingdom into the hand of Absalom thy son: and, behold, thou art taken in thy mischief, because thou art a bloody man.

mus Saul quoniam inuasisti regnum pro eo et dedit Dominus regnum in manu Absalom filii tui et ecce premunt te mala

16:9 Then said Abishai the son of Zeruiah unto the king, tua quoniam uir sanguinum es 16:9 dixit autem Abisai fi-Why should this dead dog curse my lord the king? let me go over, I pray thee, and take off his head.

lius Saruiae regi quare maledicit canis hic moriturus domino

16:10 And the king said, What have I to do with you, ye meo regi uadam et amputabo caput eius 16:10 et ait rex quid

sons of Zeruiah? so let him curse, because the LORD hath said unto him, Curse David. Who shall then say, Where-

fore hast thou done so? mihi et uobis filii Saruiae dimittite eum maledicat Dominus

enim praecepit ei ut malediceret Dauid et quis est qui audeat

16:11 And David said to Abishai, and to all his servants, dicere quare sic fecerit 16:11 et ait rex Abisai et uniuersis

Behold, my son, which came forth of my bowels, seeketh my life: how much more now may this Benjamite do it? let him alone, and let him curse; for the LORD hath bidden him.

Liber II Samuelis

seruis suis ecce filius meus qui egressus est de utero meo quaerit animam meam quanto magis nunc filius Iemini dimittite eum ut maledicat iuxta praeceptum Domini 16:12 si 16:12 It may be that the LORD will look on mine afflic-

tion, and that the LORD will requite me good for his cursing this day.

forte respiciat Dominus adflictionem meam et reddat mihi bonum pro maledictione hac hodierna 16:13 ambulabat ita- 16:13 And as David and his men went by the way, Shimei

went along on the hill’s side over against him, and cursed as he went, and threw stones at him, and cast dust.

que Dauid et socii eius per uiam cum eo Semei autem per iugum montis ex latere contra illum gradiebatur maledicens et mittens lapides aduersum eum terramque spargens 16:14 16:14 And the king, and all the people that were with him,

came weary, and refreshed themselves there.

uenit itaque rex et uniuersus populus cum eo lassus et refocilati sunt ibi 16:15 Absalom autem et omnis populus Israhel 16:15 And Absalom, and all the people the men of Israel,

came to Jerusalem, and Ahithophel with him.

ingressi sunt Hierusalem sed et Ahitofel cum eo 16:16 cum 16:16 And it came to pass, when Hushai the Archite,

David’s friend, was come unto Absalom, that Hushai said unto Absalom, God save the king, God save the king.

autem uenisset Husai Arachites amicus Dauid ad Absalom locutus est ad eum salue rex salue rex 16:17 ad quem Absa- 16:17 And Absalom said to Hushai, Is this thy kindness to

thy friend? why wentest thou not with thy friend?

lom haec est inquit gratia tua ad amicum tuum quare non isti cum amico tuo 16:18 responditque Husai ad Absalom ne- 16:18 And Hushai said unto Absalom, Nay; but whom the

LORD, and this people, and all the men of Israel, choose, his will I be, and with him will I abide.

quaquam quia illius ero quem elegit Dominus et omnis hic populus et uniuersus Israhel et cum eo manebo 16:19 sed ut 16:19 And again, whom should I serve? should I not serve

in the presence of his son? as I have served in thy father’s presence, so will I be in thy presence.

et hoc inferam cui ego seruiturus sum nonne filio regis sicut parui patri tuo sic parebo et tibi 16:20 dixit autem Absalom 16:20 Then said Absalom to Ahithophel, Give counsel

among you what we shall do.

ad Ahitofel inite consilium quid agere debeamus 16:21 et 16:21 And Ahithophel said unto Absalom, Go in unto thy

father’s concubines, which he hath left to keep the house; and all Israel shall hear that thou art abhorred of thy father:

ait Ahitofel ad Absalom ingredere ad concubinas patris tui then shall the hands of all that are with thee be strong. quas dimisit ad custodiendam domum ut cum audierit omnis Israhel quod foedaueris patrem tuum roborentur manus eorum tecum 16:22 tetenderunt igitur Absalom tabernacu- 16:22 So they spread Absalom a tent upon the top of the

house; and Absalom went in unto his father’s concubines in the sight of all Israel.

lum in solario ingressusque est ad concubinas patris sui coram uniuerso Israhel 16:23 consilium autem Ahitofel quod 16:23 And the counsel of Ahithophel, which he counselled

in those days, was as if a man had enquired at the oracle of God: so was all the counsel of Ahithophel both with

dabat in diebus illis quasi si quis consuleret Deum sic erat David and with Absalom. omne consilium Ahitofel et cum esset cum Dauid et cum esset cum Absalom

17:1 dixit igitur Ahitofel ad Absalom eligam mihi duode- 17:1 Moreover Ahithophel said unto Absalom, Let me

now choose out twelve thousand men, and I will arise and pursue after David this night:

cim milia uirorum et consurgens persequar Dauid hac nocte 17:2 et inruens super eum quippe qui lassus est et solutis ma- 17:2 And I will come upon him while he is weary and

weak handed, and will make him afraid: and all the people that are with him shall flee; and I will smite the king only:

nibus percutiam eum cumque fugerit omnis populus qui cum

Biblia Sacra Vulgata

17:3 And I will bring back all the people unto thee: the eo est percutiam regem desolatum 17:3 et reducam uniuer-

man whom thou seekest is as if all returned: so all the people shall be in peace.

sum populum quomodo omnis reuerti solet unum enim ui-

17:4 And the saying pleased Absalom well, and all the rum tu quaeris et omnis populus erit in pace 17:4 placuit-

elders of Israel.

que sermo eius Absalom et cunctis maioribus natu Israhel

17:5 Then said Absalom, Call now Hushai the Archite 17:5 ait autem Absalom uocate et Husai Arachiten et audia-

also, and let us hear likewise what he saith.

17:6 And when Hushai was come to Absalom, Absalom mus quid etiam ipse dicat 17:6 cumque uenisset Husai ad

spake unto him, saying, Ahithophel hath spoken after this manner: shall we do after his saying? if not; speak thou.

Absalom ait Absalom ad eum huiuscemodi sermonem locutus est Ahitofel facere debemus an non quod das consilium

17:7 And Hushai said unto Absalom, The counsel that 17:7 et dixit Husai ad Absalom non bonum consilium quod

Ahithophel hath given is not good at this time.

17:8 For, said Hushai, thou knowest thy father and his dedit Ahitofel hac uice 17:8 et rursum intulit Husai tu nosti

men, that they be mighty men, and they be chafed in their minds, as a bear robbed of her whelps in the field: and thy father is a man of war, and will not lodge with the people. patrem tuum et uiros qui cum eo sunt esse fortissimos et

amaro animo ueluti si ursa raptis catulis in saltu saeuiat sed

17:9 Behold, he is hid now in some pit, or in some other et pater tuus uir bellator est nec morabitur cum populo 17:9

place: and it will come to pass, when some of them be overthrown at the first, that whosoever heareth it will say, There is a slaughter among the people that follow Absalom.

forsitan nunc latitat in foueis aut in uno quo uoluerit loco et cum ceciderit unus quilibet in principio audiet quicumque audierit et dicet facta est plaga in populo qui sequebatur Ab-

17:10 And he also that is valiant, whose heart is as the salom 17:10 et fortissimus quoque cuius cor est quasi leonis

heart of a lion, shall utterly melt: for all Israel knoweth that thy father is a mighty man, and they which be with

him are valiant men. pauore soluetur scit enim omnis populus Israhel fortem esse

17:11 Therefore I counsel that all Israel be generally gath- patrem tuum et robustos omnes qui cum eo sunt 17:11 sed ered unto thee, from Dan even to Beersheba, as the sand that is by the sea for multitude; and that thou go to battle

in thine own person. hoc mihi uidetur rectum esse consilium congregetur ad te

uniuersus Israhel a Dan usque Bersabee quasi harena maris

17:12 So shall we come upon him in some place where innumerabilis et tu eris in medio eorum 17:12 et inruemus

he shall be found, and we will light upon him as the dew falleth on the ground: and of him and of all the men that

are with him there shall not be left so much as one. super eum in quocumque loco fuerit inuentus et operiemus

eum sicut cadere solet ros super terram et non relinquemus

17:13 Moreover, if he be gotten into a city, then shall all de uiris qui cum eo sunt ne unum quidem 17:13 quod si Israel bring ropes to that city, and we will draw it into the river, until there be not one small stone found there.

urbem aliquam fuerit ingressus circumdabit omnis Israhel ciuitati illi funes et trahemus eam in torrentem ut non rep-

17:14 And Absalom and all the men of Israel said, The periatur nec calculus quidem ex ea 17:14 dixitque Absalom

counsel of Hushai the Archite is better than the counsel of Ahithophel. For the LORD had appointed to defeat the good counsel of Ahithophel, to the intent that the LORD might bring evil upon Absalom.

et omnis uir Israhel melius consilium Husai Arachitae consilio Ahitofel Domini autem nutu dissipatum est consilium Ahitofel utile ut induceret Dominus super Absalom malum

17:15 Then said Hushai unto Zadok and to Abiathar the 17:15 et ait Husai Sadoc et Abiathar sacerdotibus hoc et hoc

priests, Thus and thus did Ahithophel counsel Absalom and the elders of Israel; and thus and thus have I coun-

selled. modo consilium dedit Ahitofel Absalom et senibus Israhel et

Liber II Samuelis

ego tale et tale dedi consilium 17:16 nunc ergo mittite cito 17:16 Now therefore send quickly, and tell David, saying,

Lodge not this night in the plains of the wilderness, but speedily pass over; lest the king be swallowed up, and all

et nuntiate Dauid dicentes ne moremini nocte hac in cam- the people that are with him. pestribus deserti sed absque dilatione transgredere ne forte absorbeatur rex et omnis populus qui cum eo est 17:17 Io- 17:17 Now Jonathan and Ahimaaz stayed by Enrogel; for

they might not be seen to come into the city: and a wench went and told them; and they went and told king David.

nathan autem et Achimaas stabant iuxta fontem Rogel abiit ancilla et nuntiauit eis et illi profecti sunt ut referrent ad regem Dauid nuntium non enim poterant uideri aut introire ciuitatem 17:18 uidit autem eos quidam puer et indicauit 17:18 Nevertheless a lad saw them, and told Absalom: but

they went both of them away quickly, and came to a man’s house in Bahurim, which had a well in his court; whither

Absalom illi uero concito gradu ingressi sunt domum cuius- they went down. dam uiri in Baurim qui habebat puteum in uestibulo suo et descenderunt in eum 17:19 tulit autem mulier et expandit 17:19 And the woman took and spread a covering over

the well’s mouth, and spread ground corn thereon; and the thing was not known.

uelamen super os putei quasi siccans ptisanas et sic res latuit 17:20 cumque uenissent serui Absalom ad mulierem in 17:20 And when Absalom’s servants came to the woman

to the house, they said, Where is Ahimaaz and Jonathan? And the woman said unto them, They be gone over the brook of water. And when they had sought and could not find them, they returned to Jerusalem.

domum dixerunt ubi est Achimaas et Ionathan et respondit eis mulier transierunt gustata paululum aqua at hii qui quaerebant cum non repperissent reuersi sunt Hierusalem 17:21 17:21 And it came to pass, after they were departed, that

they came up out of the well, and went and told king David, and said unto David, Arise, and pass quickly over the water: for thus hath Ahithophel counselled against you.

cumque abissent ascenderunt illi de puteo et pergentes nuntiauerunt regi Dauid atque dixerunt surgite transite cito fluuium quoniam huiuscemodi dedit consilium contra uos Ahitofel 17:22 surrexit ergo Dauid et omnis populus qui erat 17:22 Then David arose, and all the people that were with

him, and they passed over Jordan: by the morning light there lacked not one of them that was not gone over Jor-

cum eo et transierunt Iordanem donec dilucesceret et ne unus dan. quidem residuus fuit qui non transisset fluuium 17:23 porro 17:23 And when Ahithophel saw that his counsel was not

followed, he saddled his ass, and arose, and gat him home to his house, to his city, and put his household in order, and hanged himself, and died, and was buried in the sepulchre of his father.

Ahitofel uidens quod non fuisset factum consilium suum strauit asinum suum et surrexit et abiit in domum suam et in ciuitatem suam et disposita domo sua suspendio interiit et sepultus est in sepulchro patris sui 17:24 Dauid autem ue- 17:24 Then David came to Mahanaim. And Absalom

passed over Jordan, he and all the men of Israel with him.

nit in Castra et Absalom transiuit Iordanem ipse et omnis uir Israhel cum eo 17:25 Amasam uero constituit Absalom 17:25 And Absalom made Amasa captain of the host in-

stead of Joab: which Amasa was a man’s son, whose name was Ithra an Israelite, that went in to Abigail the daughter

pro Ioab super exercitum Amasa autem erat filius uiri qui of Nahash, sister to Zeruiah Joab’s mother. uocabatur Iethra de Hiesreli qui ingressus est ad Abigail filiam Naas sororem Saruiae quae fuit mater Ioab 17:26 et ca- 17:26 So Israel and Absalom pitched in the land of Gilead. strametatus est Israhel cum Absalom in terra Galaad 17:27 17:27 And it came to pass, when David was come to Ma-

hanaim, that Shobi the son of Nahash of Rabbah of the children of Ammon, and Machir the son of Ammiel of

cumque uenisset Dauid in Castra Sobi filius Naas de Rab- Lodebar, and Barzillai the Gileadite of Rogelim,

Biblia Sacra Vulgata

bath filiorum Ammon et Machir filius Ammihel de Lodabar

17:28 Brought beds, and basons, and earthen vessels, and et Berzellai Galaadites de Rogelim 17:28 obtulerunt ei st-

wheat, and barley, and flour, and parched corn, and beans, and lentiles, and parched pulse,

ratoria et tappetia et uasa fictilia frumentum et hordeum et

17:29 And honey, and butter, and sheep, and cheese of farinam pulentam et fabam et lentem frixum cicer 17:29 et

kine, for David, and for the people that were with him, to eat: for they said, The people is hungry, and weary, and

thirsty, in the wilderness. mel et butyrum oues et pingues uitulos dederuntque Dauid et populo qui cum eo erat ad uescendum suspicati enim sunt populum fame et siti fatigari in deserto

18:1 And David numbered the people that were with him, 18:1 igitur considerato Dauid populo suo constituit super

and set captains of thousands, and captains of hundreds over them.

eum tribunos et centuriones 18:2 et dedit populi tertiam

18:2 And David sent forth a third part of the people under the hand of Joab, and a third part under the hand of Abishai the son of Zeruiah, Joab’s brother, and a third part under the hand of Ittai the Gittite. And the king said unto the people, I will surely go forth with you myself also.

partem sub manu Ioab et tertiam in manu Abisai filii Saruiae fratris Ioab et tertiam sub manu Ethai qui erat de Geth

18:3 But the people answered, Thou shalt not go forth: for dixitque rex ad populum egrediar et ego uobiscum 18:3 et if we flee away, they will not care for us; neither if half of us die, will they care for us: but now thou art worth ten thousand of us: therefore now it is better that thou succour us out of the city.

respondit populus non exibis siue enim fugerimus non magnopere ad eos de nobis pertinebit siue media pars ceciderit e nobis non satis curabunt quia tu unus pro decem milibus con-

18:4 And the king said unto them, What seemeth you best putaris melius est igitur ut sis nobis in urbe praesidio 18:4

I will do. And the king stood by the gate side, and all the people came out by hundreds and by thousands.

ad quos rex ait quod uobis rectum uidetur hoc faciam stetit ergo rex iuxta portam egrediebaturque populus per turmas

18:5 And the king commanded Joab and Abishai and Ittai, suas centeni et milleni 18:5 et praecepit rex Ioab et Abisai

saying, Deal gently for my sake with the young man, even with Absalom. And all the people heard when the king

gave all the captains charge concerning Absalom. et Ethai dicens seruate mihi puerum Absalom et omnis po-

pulus audiebat praecipientem regem cunctis principibus pro

18:6 So the people went out into the field against Israel: Absalom 18:6 itaque egressus est populus in campum conand the battle was in the wood of Ephraim;

18:7 Where the people of Israel were slain before the ser- tra Israhel et factum est proelium in saltu Ephraim 18:7 et

vants of David, and there was there a great slaughter that day of twenty thousand men.

caesus est ibi populus Israhel ab exercitu Dauid factaque est

18:8 For the battle was there scattered over the face of all ibi plaga magna in die illa uiginti milium 18:8 fuit autem ibi

the country: and the wood devoured more people that day than the sword devoured.

proelium dispersum super faciem omnis terrae et multo plures erant quos saltus consumpserat de populo quam hii quos

18:9 And Absalom met the servants of David. And Absa- uorauerat gladius in die illa 18:9 accidit autem ut occur-

lom rode upon a mule, and the mule went under the thick boughs of a great oak, and his head caught hold of the oak, and he was taken up between the heaven and the earth; and the mule that was under him went away.

reret Absalom seruis Dauid sedens mulo cumque ingressus fuisset mulus subter condensam quercum et magnam adhesit caput eius quercui et illo suspenso inter caelum et ter-

18:10 And a certain man saw it, and told Joab, and said, ram mulus cui sederat pertransiuit 18:10 uidit autem hoc

Behold, I saw Absalom hanged in an oak.

quispiam et nuntiauit Ioab dicens uidi Absalom pendere de

Liber II Samuelis

quercu 18:11 et ait Ioab uiro qui nuntiauerat ei si uidisti 18:11 And Joab said unto the man that told him, And,

behold, thou sawest him, and why didst thou not smite him there to the ground? and I would have given thee ten

quare non confodisti eum cum terra et ego dedissem tibi de- shekels of silver, and a girdle. cem argenti siclos et unum balteum 18:12 qui dixit ad Ioab 18:12 And the man said unto Joab, Though I should re-

ceive a thousand shekels of silver in mine hand, yet would I not put forth mine hand against the king’s son: for in our hearing the king charged thee and Abishai and Ittai, saying, Beware that none touch the young man Absalom.

si adpenderes in manibus meis mille argenteos nequaquam mitterem manum meam in filium regis audientibus enim nobis praecepit rex tibi et Abisai et Ethai dicens custodite mihi

puerum Absalom 18:13 sed et si fecissem contra animam 18:13 Otherwise I should have wrought falsehood against mine own life: for there is no matter hid from the king, and thou thyself wouldest have set thyself against me.

meam audacter nequaquam hoc regem latere potuisset et tu stares ex aduerso 18:14 et ait Ioab non sicut tu uis sed ad- 18:14 Then said Joab, I may not tarry thus with thee. And

he took three darts in his hand, and thrust them through the heart of Absalom, while he was yet alive in the midst

grediar eum coram te tulit ergo tres lanceas in manu sua et of the oak. infixit eas in corde Absalom cumque adhuc palpitaret herens in quercu 18:15 cucurrerunt decem iuuenes armigeri Ioab 18:15 And ten young men that bare Joab’s armour com-

passed about and smote Absalom, and slew him.

et percutientes interfecerunt eum 18:16 cecinit autem Ioab 18:16 And Joab blew the trumpet, and the people returned

from pursuing after Israel: for Joab held back the people.

bucina et retinuit populum ne persequeretur fugientem Israhel uolens parcere multitudini 18:17 et tulerunt Absalom 18:17 And they took Absalom, and cast him into a great

pit in the wood, and laid a very great heap of stones upon him: and all Israel fled every one to his tent.

et proiecerunt eum in saltu in foueam grandem et conportauerunt super eum aceruum lapidum magnum nimis omnis autem Israhel fugit in tabernacula sua 18:18 porro Absalom 18:18 Now Absalom in his lifetime had taken and reared

up for himself a pillar, which is in the king’s dale: for he said, I have no son to keep my name in remembrance: and he called the pillar after his own name: and it is called unto this day, Absalom’s place.

erexerat sibi cum adhuc uiueret titulum qui est in ualle Regis dixerat enim non habeo filium et hoc erit monumentum nominis mei uocauitque titulum nomine suo et appellatur manus Absalom usque ad hanc diem 18:19 Achimaas autem 18:19 Then said Ahimaaz the son of Zadok, Let me now

run, and bear the king tidings, how that the LORD hath avenged him of his enemies.

filius Sadoc ait curram et nuntiabo regi quia iudicium fecerit

ei Dominus de manu inimicorum eius 18:20 ad quem Ioab 18:20 And Joab said unto him, Thou shalt not bear tidings this day, but thou shalt bear tidings another day: but this day thou shalt bear no tidings, because the king’s son is

dixit non eris nuntius in hac die sed nuntiabis in alia hodie dead. nolo te nuntiare filius enim regis est mortuus 18:21 et ait 18:21 Then said Joab to Cushi, Go tell the king what thou

hast seen. And Cushi bowed himself unto Joab, and ran.

Ioab Chusi uade et nuntia regi quae uidisti adorauit Chusi Ioab et cucurrit 18:22 rursum autem Achimaas filius Sadoc 18:22 Then said Ahimaaz the son of Zadok yet again to

Joab, But howsoever, let me, I pray thee, also run after Cushi. And Joab said, Wherefore wilt thou run, my son,

dixit ad Ioab quid inpedit si etiam ego curram post Chusi seeing that thou hast no tidings ready? dixitque Ioab quid uis currere fili mi non eris boni nuntii baiulus 18:23 qui respondit quid enim si cucurrero et ait ei 18:23 But howsoever, said he, let me run. And he said unto

him, Run. Then Ahimaaz ran by the way of the plain, and overran Cushi.

curre currens ergo Achimaas per uiam conpendii transiuit Chusi 18:24 Dauid autem sedebat inter duas portas specu- 18:24 And David sat between the two gates: and the

watchman went up to the roof over the gate unto the wall, and lifted up his eyes, and looked, and behold a man running alone.

Biblia Sacra Vulgata

lator uero qui erat in fastigio portae super murum eleuans

18:25 And the watchman cried, and told the king. And the oculos uidit hominem currentem solum 18:25 et exclamans

king said, If he be alone, there is tidings in his mouth. And he came apace, and drew near.

indicauit regi dixitque rex si solus est bonus est nuntius in

18:26 And the watchman saw another man running: and ore eius properante autem illo et accedente propius 18:26

the watchman called unto the porter, and said, Behold another man running alone. And the king said, He also

bringeth tidings. uidit speculator hominem alterum currentem et uociferans

in culmine ait apparet mihi homo currens solus dixitque rex

18:27 And the watchman said, Me thinketh the running et iste bonus est nuntius 18:27 speculator autem contemplor

of the foremost is like the running of Ahimaaz the son of Zadok. And the king said, He is a good man, and cometh

with good tidings. ait cursum prioris quasi cursum Achimaas filii Sadoc et ait 18:28 And Ahimaaz called, and said unto the king, All is rex uir bonus est et nuntium portans bonum uenit 18:28 cla-

well. And he fell down to the earth upon his face before the king, and said, Blessed be the LORD thy God, which hath delivered up the men that lifted up their hand against my lord the king.

mans autem Achimaas dixit ad regem salue et adorans regem coram eo pronus in terram ait benedictus Dominus Deus tuus qui conclusit homines qui leuauerunt manus suas contra do-

18:29 And the king said, Is the young man Absalom safe? minum meum regem 18:29 et ait rex estne pax puero Ab-And Ahimaaz answered, When Joab sent the king’s servant, and me thy servant, I saw a great tumult, but I knew

not what it was. salom dixitque Achimaas uidi tumultum magnum cum mit-

teret Ioab seruus tuus o rex me seruum tuum nescio aliud

18:30 And the king said unto him, Turn aside, and stand 18:30 ad quem rex transi ait et sta hic cumque ille transisset

here. And he turned aside, and stood still.

18:31 And, behold, Cushi came; and Cushi said, Tidings, et staret 18:31 apparuit Chusi et ueniens ait bonum adporto

my lord the king: for the LORD hath avenged thee this day of all them that rose up against thee.

nuntium domine mi rex iudicauit enim pro te Dominus ho-

18:32 And the king said unto Cushi, Is the young man die de manu omnium qui surrexerunt contra te 18:32 dixit

Absalom safe? And Cushi answered, The enemies of my lord the king, and all that rise against thee to do thee hurt,

be as that young man is. autem rex ad Chusi estne pax puero Absalom cui respon-

dens Chusi fiant inquit sicut puer inimici domini mei regis

18:33 And the king was much moved, and went up to the et uniuersi qui consurgunt aduersum eum in malum 18:33

chamber over the gate, and wept: and as he went, thus he said, O my son Absalom, my son, my son Absalom! would God I had died for thee, O Absalom, my son, my son!

contristatus itaque rex ascendit cenaculum portae et fleuit et sic loquebatur uadens fili mi Absalom fili mi Absalom quis mihi tribuat ut ego moriar pro te Absalom fili mi fili mi

19:1 And it was told Joab, Behold, the king weepeth and 19:1 nuntiatum est autem Ioab quod rex fleret et lugeret

mourneth for Absalom.

19:2 And the victory that day was turned into mourning filium suum 19:2 et uersa est uictoria in die illa in luctum

unto all the people: for the people heard say that day how the king was grieved for his son.

omni populo audiuit enim populus in die illa dici dolet rex

19:3 And the people gat them by stealth that day into the super filio suo 19:3 et declinabat populus in die illa ingredi

city, as people being ashamed steal away when they flee in battle.

ciuitatem quomodo declinare solet populus uersus et fugiens

19:4 But the king covered his face, and the king cried with de proelio 19:4 porro rex operuit caput suum et clamabat a loud voice, O my son Absalom, O Absalom, my son, my son!

uoce magna fili mi Absalom Absalom fili mi fili mi 19:5

19:5 And Joab came into the house to the king, and said, Thou hast shamed this day the faces of all thy servants, which this day have saved thy life, and the lives of thy sons and of thy daughters, and the lives of thy wives, and the lives of thy concubines;

ingressus ergo Ioab ad regem in domo dixit confudisti hodie

Liber II Samuelis

uultus omnium seruorum tuorum qui saluam fecerunt animam tuam et animam filiorum tuorum et filiarum tuarum et animam uxorum tuarum et animam concubinarum tuarum 19:6 diligis odientes te et odio habes diligentes te et osten- 19:6 In that thou lovest thine enemies, and hatest thy

friends. For thou hast declared this day, that thou regardest neither princes nor servants: for this day I perceive, that if Absalom had lived, and all we had died this day, then it had pleased thee well.

disti hodie quia non curas de ducibus tuis et de seruis tuis et uere cognoui modo quia si Absalom uiueret et nos omnes occubuissemus tunc placeret tibi 19:7 nunc igitur surge et 19:7 Now therefore arise, go forth, and speak comfortably

unto thy servants: for I swear by the LORD, if thou go not forth, there will not tarry one with thee this night: and that will be worse unto thee than all the evil that befell thee from thy youth until now.

procede et adloquens satisfac seruis tuis iuro enim tibi per Dominum quod si non exieris ne unus quidem remansurus sit tecum nocte hac et peius erit hoc tibi quam omnia mala quae uenerunt super te ab adulescentia tua usque in praesens 19:8 surrexit ergo rex et sedit in porta et omni populo 19:8 Then the king arose, and sat in the gate. And they

told unto all the people, saying, Behold, the king doth sit in the gate. And all the people came before the king: for

nuntiatum est quod rex sederet in porta uenitque uniuersa Israel had fled every man to his tent. multitudo coram rege Israhel autem fugit in tabernacula sua 19:9 omnis quoque populus certabat in cunctis tribubus Isra- 19:9 And all the people were at strife throughout all the

tribes of Israel, saying, The king saved us out of the hand of our enemies, and he delivered us out of the hand of the

hel dicens rex liberauit nos de manu inimicorum nostrorum Philistines; and now he is fled out of the land for Absalom. ipse saluauit nos de manu Philisthinorum et nunc fugit de terra propter Absalom 19:10 Absalom autem quem unxi- 19:10 And Absalom, whom we anointed over us, is dead in

battle. Now therefore why speak ye not a word of bringing the king back?

mus super nos mortuus est in bello usquequo siletis et non reducitis regem 19:11 rex uero Dauid misit ad Sadoc et ad 19:11 And king David sent to Zadok and to Abiathar the

priests, saying, Speak unto the elders of Judah, saying, Why are ye the last to bring the king back to his house? seeing the speech of all Israel is come to the king, even to his house.

Abiathar sacerdotes dicens loquimini ad maiores natu Iuda dicentes cur uenitis nouissimi ad reducendum regem in domum suam sermo autem omnis Israhel peruenerat ad regem in domo eius 19:12 fratres mei uos os meum et caro mea 19:12 Ye are my brethren, ye are my bones and my flesh:

wherefore then are ye the last to bring back the king?

uos quare nouissimi reducitis regem 19:13 et Amasae di- 19:13 And say ye to Amasa, Art thou not of my bone, and

of my flesh? God do so to me, and more also, if thou be not captain of the host before me continually in the room

cite nonne os meum es et caro mea haec faciat mihi Deus of Joab. et haec addat si non magister militiae fueris coram me omni tempore pro Ioab 19:14 et inclinauit cor omnium uirorum 19:14 And he bowed the heart of all the men of Judah,

even as the heart of one man; so that they sent this word unto the king, Return thou, and all thy servants.

Iuda quasi uiri unius miseruntque ad regem dicentes reuertere tu et omnes serui tui 19:15 et reuersus est rex et ue- 19:15 So the king returned, and came to Jordan. And Ju-

dah came to Gilgal, to go to meet the king, to conduct the king over Jordan.

nit usque ad Iordanem et Iuda uenit in Galgala ut occurreret regi et transduceret eum Iordanem 19:16 festinauit autem 19:16 And Shimei the son of Gera, a Benjamite, which

was of Bahurim, hasted and came down with the men of Judah to meet king David.

Semei filius Gera filii Iemini de Baurim et descendit cum ui-

Biblia Sacra Vulgata

19:17 And there were a thousand men of Benjamin with ris Iuda in occursum regis Dauid 19:17 cum mille uiris de

him, and Ziba the servant of the house of Saul, and his fifteen sons and his twenty servants with him; and they

went over Jordan before the king. Beniamin et Siba puer de domo Saul et quindecim filii eius

ac uiginti serui erant cum eo et inrumpentes Iordanem ante

19:18 And there went over a ferry boat to carry over the regem 19:18 transierunt uada ut transducerent domum regis

king’s household, and to do what he thought good. And Shimei the son of Gera fell down before the king, as he

was come over Jordan; et facerent iuxta iussionem eius Semei autem filius Gera pro-19:19 And said unto the king, Let not my lord impute in- stratus coram rege cum iam transisset Iordanem 19:19 dixit

iquity unto me, neither do thou remember that which thy servant did perversely the day that my lord the king went out of Jerusalem, that the king should take it to his heart. ad eum ne reputes mihi domine mi iniquitatem neque memi-

neris iniuriam serui tui in die qua egressus es domine mi rex

19:20 For thy servant doth know that I have sinned: there- de Hierusalem neque ponas rex in corde tuo 19:20 agnosco fore, behold, I am come the first this day of all the house of Joseph to go down to meet my lord the king.

enim seruus tuus peccatum meum et idcirco hodie primus ueni de omni domo Ioseph descendique in occursum domini

19:21 But Abishai the son of Zeruiah answered and said, mei regis 19:21 respondens uero Abisai filius Saruiae dixit

Shall not Shimei be put to death for this, because he cursed the LORD’s anointed?

numquid pro his uerbis non occidetur Semei quia maledixit

19:22 And David said, What have I to do with you, ye sons christo Domini 19:22 et ait Dauid quid mihi et uobis filii

of Zeruiah, that ye should this day be adversaries unto me? shall there any man be put to death this day in Israel? for

do not I know that I am this day king over Israel? Saruiae cur efficimini mihi hodie in Satan ergone hodie in-

terficietur uir in Israhel an ignoro hodie me factum regem

19:23 Therefore the king said unto Shimei, Thou shalt not super Israhel 19:23 et ait rex Semei non morieris iurauitque

die. And the king sware unto him.

19:24 And Mephibosheth the son of Saul came down ei 19:24 Mifiboseth quoque filius Saul descendit in occur-

to meet the king, and had neither dressed his feet, nor trimmed his beard, nor washed his clothes, from the day

the king departed until the day he came again in peace. sum regis inlotis pedibus et intonsa barba uestesque suas non

lauerat a die qua egressus fuerat rex usque ad diem reuersio-

19:25 And it came to pass, when he was come to Jerusalem nis eius in pace 19:25 cumque Hierusalem occurrisset regi

to meet the king, that the king said unto him, Wherefore wentest not thou with me, Mephibosheth?

dixit ei rex quare non uenisti mecum Mifiboseth 19:26 qui

19:26 And he answered, My lord, O king, my servant deceived me: for thy servant said, I will saddle me an ass, that I may ride thereon, and go to the king; because thy servant is lame.

respondens ait domine mi rex seruus meus contempsit me dixi ei ego famulus tuus ut sterneret mihi asinum et ascen-

19:27 And he hath slandered thy servant unto my lord the dens abirem cum rege claudus enim sum seruus tuus 19:27 king; but my lord the king is as an angel of God: do therefore what is good in thine eyes.

insuper et accusauit me seruum tuum ad te dominum meum regem tu autem domine mi rex sicut angelus Dei fac quod

19:28 For all of my father’s house were but dead men be- placitum est tibi 19:28 neque enim fuit domus patris mei

fore my lord the king: yet didst thou set thy servant among them that did eat at thine own table. What right therefore

have I yet to cry any more unto the king? nisi morti obnoxia domino meo regi tu autem posuisti me seruum tuum inter conuiuas mensae tuae quid igitur habeo iustae querellae aut quid possum ultra uociferari ad regem

19:29 And the king said unto him, Why speakest thou any 19:29 ait ergo ei rex quid ultra loqueris fixum est quod locu-

more of thy matters? I have said, Thou and Ziba divide the land.

tus sum tu et Siba diuidite possessiones 19:30 responditque

19:30 And Mephibosheth said unto the king, Yea, let him take all, forasmuch as my lord the king is come again in peace unto his own house.

Liber II Samuelis

Mifiboseth regi etiam cuncta accipiat postquam reuersus est dominus meus rex pacifice in domum suam 19:31 Berzellai 19:31 And Barzillai the Gileadite came down from Roge-

lim, and went over Jordan with the king, to conduct him over Jordan.

quoque Galaadites descendens de Rogelim transduxit regem Iordanem paratus etiam ultra fluuium prosequi eum 19:32 19:32 Now Barzillai was a very aged man, even fourscore

years old: and he had provided the king of sustenance while he lay at Mahanaim; for he was a very great man.

erat autem Berzellai Galaadites senex ualde id est octogenarius et ipse praebuit alimenta regi cum moraretur in Castris fuit quippe uir diues nimis 19:33 dixit itaque rex ad Berzel- 19:33 And the king said unto Barzillai, Come thou over

with me, and I will feed thee with me in Jerusalem.

lai ueni mecum ut requiescas secure mecum in Hierusalem 19:34 et ait Berzellai ad regem quot sunt dies annorum uitae 19:34 And Barzillai said unto the king, How long have I

to live, that I should go up with the king unto Jerusalem?

meae ut ascendam cum rege Hierusalem 19:35 octogena- 19:35 I am this day fourscore years old: and can I discern between good and evil? can thy servant taste what I eat or what I drink? can I hear any more the voice of singing men and singing women? wherefore then should thy servant be yet a burden unto my lord the king?

rius sum hodie numquid uigent sensus mei ad discernendum suaue aut amarum aut delectare potest seruum tuum cibus et potus uel audire ultra possum uocem cantorum atque cantricum quare seruus tuus fit oneri domino meo regi 19:36 pau- 19:36 Thy servant will go a little way over Jordan with

the king: and why should the king recompense it me with such a reward?

lulum procedam famulus tuus ab Iordane tecum nec indigeo hac uicissitudine 19:37 sed obsecro ut reuertar seruus tuus 19:37 Let thy servant, I pray thee, turn back again, that

I may die in mine own city, and be buried by the grave of my father and of my mother. But behold thy servant Chimham; let him go over with my lord the king; and do to him what shall seem good unto thee.

et moriar in ciuitate mea iuxta sepulchrum patris mei et matris meae est autem seruus tuus Chamaam ipse uadat tecum domine mi rex et fac ei quod tibi bonum uidetur 19:38 dixit- 19:38 And the king answered, Chimham shall go over with

me, and I will do to him that which shall seem good unto thee: and whatsoever thou shalt require of me, that will I

que rex mecum transeat Chamaam et ego faciam ei quicquid do for thee. tibi placuerit et omne quod petieris a me inpetrabis 19:39 19:39 And all the people went over Jordan. And when the

king was come over, the king kissed Barzillai, and blessed him; and he returned unto his own place.

cumque transisset uniuersus populus et rex Iordanem osculatus est rex Berzellai et benedixit ei et ille reuersus est in locum suum 19:40 transiuit ergo rex in Galgalam et Cha- 19:40 Then the king went on to Gilgal, and Chimham went

on with him: and all the people of Judah conducted the king, and also half the people of Israel.

maam cum eo omnis autem populus Iuda transduxerat regem et media tantum pars adfuerat de populo Israhel 19:41 19:41 And, behold, all the men of Israel came to the king,

and said unto the king, Why have our brethren the men of Judah stolen thee away, and have brought the king, and his

itaque omnes uiri Israhel concurrentes ad regem dixerunt ei household, and all David’s men with him, over Jordan? quare te furati sunt fratres nostri uiri Iuda et transduxerunt regem et domum eius Iordanem omnesque uiros Dauid cum eo 19:42 et respondit omnis uir Iuda ad uiros Israhel quia 19:42 And all the men of Judah answered the men of Is-

rael, Because the king is near of kin to us: wherefore then be ye angry for this matter? have we eaten at all of the

propior mihi est rex cur irasceris super hac re numquid co- king’s cost? or hath he given us any gift? medimus aliquid ex rege aut munera nobis data sunt 19:43 19:43 And the men of Israel answered the men of Judah,

and said, We have ten parts in the king, and we have also more right in David than ye: why then did ye despise us, that our advice should not be first had in bringing back our king? And the words of the men of Judah were fiercer than the words of the men of Israel.

et respondit uir Israhel ad uiros Iuda et ait decem partibus

Biblia Sacra Vulgata

maior ego sum apud regem magisque ad me pertinet Dauid quam ad te cur mihi fecisti iniuriam et non mihi nuntiatum est priori ut reducerem regem meum durius autem responderunt uiri Iuda uiris Israhel

20:1 And there happened to be there a man of Belial, 20:1 accidit quoque ut ibi esset uir Belial nomine Seba fi-

whose name was Sheba, the son of Bichri, a Benjamite: and he blew a trumpet, and said, We have no part in David, neither have we inheritance in the son of Jesse: every man to his tents, O Israel.

lius Bochri uir iemineus et cecinit bucina et ait non est nobis pars in Dauid neque hereditas in filio Isai uir in tabernacula

20:2 So every man of Israel went up from after David, and tua Israhel 20:2 et separatus est omnis Israhel a Dauid se-

followed Sheba the son of Bichri: but the men of Judah clave unto their king, from Jordan even to Jerusalem.

cutusque est Seba filium Bochri uiri autem Iuda adheserunt

20:3 And David came to his house at Jerusalem; and the regi suo a Iordane usque Hierusalem 20:3 cumque uenisset

king took the ten women his concubines, whom he had left to keep the house, and put them in ward, and fed them, but went not in unto them. So they were shut up unto the day of their death, living in widowhood.

rex in domum suam Hierusalem tulit decem mulieres concubinas quas dereliquerat ad custodiendam domum et tradidit eas in custodiam alimenta eis praebens et non est ingressus ad eas sed erant clausae usque ad diem mortis suae

20:4 Then said the king to Amasa, Assemble me the men in uiduitate uiuentes 20:4 dixit autem rex Amasae conuoca

of Judah within three days, and be thou here present.

mihi omnes uiros Iuda in diem tertium et tu adesto prae-

20:5 So Amasa went to assemble the men of Judah: but sens 20:5 abiit ergo Amasa ut conuocaret Iudam et mora-

he tarried longer than the set time which he had appointed him.

tus est extra placitum quod ei constituerat 20:6 ait autem

20:6 And David said to Abishai, Now shall Sheba the son of Bichri do us more harm than did Absalom: take thou thy lord’s servants, and pursue after him, lest he get him fenced cities, and escape us.

Dauid ad Abisai nunc magis adflicturus est nos Seba filius Bochri quam Absalom tolle igitur seruos domini tui et persequere eum ne forte inueniat ciuitates munitas et effugiat

20:7 And there went out after him Joab’s men, and the nos 20:7 egressi sunt ergo cum eo uiri Ioab Cherethi quo-

Cherethites, and the Pelethites, and all the mighty men: and they went out of Jerusalem, to pursue after Sheba the

son of Bichri. que et Felethi et omnes robusti exierunt de Hierusalem ad

20:8 When they were at the great stone which is in Gibeon, persequendum Seba filium Bochri 20:8 cumque illi essent Amasa went before them. And Joab’s garment that he had put on was girded unto him, and upon it a girdle with a sword fastened upon his loins in the sheath thereof; and as he went forth it fell out.

iuxta lapidem grandem qui est in Gabaon Amasa ueniens occurrit eis porro Ioab uestitus erat tunica stricta ad mensuram habitus sui et desuper accinctus gladio dependente usque ad ilia in uagina qui fabrefactus leui motu egredi poterat

20:9 And Joab said to Amasa, Art thou in health, my et percutere 20:9 dixit itaque Ioab ad Amasa salue mi fra-

brother? And Joab took Amasa by the beard with the right hand to kiss him.

ter et tenuit manu dextra mentum Amasae quasi osculans

20:10 But Amasa took no heed to the sword that was in eum 20:10 porro Amasa non obseruauit gladium quem ha-

Joab’s hand: so he smote him therewith in the fifth rib, and shed out his bowels to the ground, and struck him not again; and he died. So Joab and Abishai his brother pursued after Sheba the son of Bichri.

bebat Ioab qui percussit eum in latere et effudit intestina eius in terram nec secundum uulnus adposuit Ioab autem et Abi-

Liber II Samuelis

sai frater eius persecuti sunt Seba filium Bochri 20:11 inte- 20:11 And one of Joab’s men stood by him, and said, He

that favoureth Joab, and he that is for David, let him go after Joab.

rea quidam uiri cum stetissent iuxta cadauer Amasae de sociis Ioab dixerunt ecce qui esse uoluit pro Ioab comes Dauid pro Ioab 20:12 Amasa autem conspersus sanguine iacebat 20:12 And Amasa wallowed in blood in the midst of the

highway. And when the man saw that all the people stood still, he removed Amasa out of the highway into the field, and cast a cloth upon him, when he saw that every one that came by him stood still.

in media uia uidit hoc quidam uir quod subsisteret omnis populus ad uidendum eum et amouit Amasam de uia in agrum operuitque eum uestimento ne subsisterent transeuntes propter eum 20:13 amoto igitur illo de uia transiebat om- 20:13 When he was removed out of the highway, all the

people went on after Joab, to pursue after Sheba the son of Bichri.

nis uir sequens Ioab ad persequendum Seba filium Bochri 20:14 porro ille transierat per omnes tribus Israhel in Abe- 20:14 And he went through all the tribes of Israel unto

Abel, and to Bethmaachah, and all the Berites: and they were gathered together, and went also after him.

lam et in Bethmacha omnesque electi congregati fuerant ad eum 20:15 uenerunt itaque et obpugnabant eum in Abela et 20:15 And they came and besieged him in Abel of Beth-

maachah, and they cast up a bank against the city, and it stood in the trench: and all the people that were with Joab

in Bethmacha et circumdederunt munitionibus ciuitatem et battered the wall, to throw it down. obsessa est urbs omnis autem turba quae erat cum Ioab moliebatur destruere muros 20:16 et exclamauit mulier sapiens 20:16 Then cried a wise woman out of the city, Hear, hear;

say, I pray you, unto Joab, Come near hither, that I may speak with thee.

de ciuitate audite audite dicite Ioab adpropinqua huc et loquar tecum 20:17 qui cum accessisset ad eam ait illi tu es 20:17 And when he was come near unto her, the woman

said, Art thou Joab? And he answered, I am he. Then she said unto him, Hear the words of thine handmaid. And he

Ioab et ille respondit ego ad quem sic locuta est audi sermo- answered, I do hear. nes ancillae tuae qui respondit audio 20:18 rursumque illa 20:18 Then she spake, saying, They were wont to speak

in old time, saying, They shall surely ask counsel at Abel: and so they ended the matter.

sermo inquit dicebatur in ueteri prouerbio qui interrogant interrogent in Abela et sic perficiebant 20:19 nonne ego sum 20:19 I am one of them that are peaceable and faithful in

Israel: thou seekest to destroy a city and a mother in Israel: why wilt thou swallow up the inheritance of the LORD?

quae respondeo ueritatem Israhel et tu quaeris subruere ciuitatem et euertere matrem in Israhel quare praecipitas hereditatem Domini 20:20 respondensque Ioab ait absit absit hoc 20:20 And Joab answered and said, Far be it, far be it from

me, that I should swallow up or destroy.

a me non praecipito neque demolior 20:21 non se sic habet 20:21 The matter is not so: but a man of mount Ephraim,

Sheba the son of Bichri by name, hath lifted up his hand against the king, even against David: deliver him only, and I will depart from the city. And the woman said unto Joab, Behold, his head shall be thrown to thee over the wall.

res sed homo de monte Ephraim Seba filius Bochri cognomine leuauit manum contra regem Dauid tradite illum solum et recedemus a ciuitate et ait mulier ad Ioab ecce caput eius mittetur ad te per murum 20:22 ingressa est ergo ad om- 20:22 Then the woman went unto all the people in her

wisdom. And they cut off the head of Sheba the son of Bichri, and cast it out to Joab. And he blew a trumpet, and they retired from the city, every man to his tent. And Joab returned to Jerusalem unto the king.

nem populum et locuta est eis sapienter qui abscisum caput Seba filii Bochri proiecerunt ad Ioab et ille cecinit tuba et recesserunt ab urbe unusquisque in tabernacula sua Ioab autem reuersus est Hierusalem ad regem 20:23 fuit ergo Ioab 20:23 Now Joab was over all the host of Israel: and Bena-

iah the son of Jehoiada was over the Cherethites and over the Pelethites:

Biblia Sacra Vulgata

super omnem exercitum Israhel Banaias autem filius Ioiadae

20:24 And Adoram was over the tribute: and Jehoshaphat super Cheretheos et Feletheos 20:24 Aduram uero super tri-

the son of Ahilud was recorder:

20:25 And Sheva was scribe: and Zadok and Abiathar buta porro Iosaphat filius Ahilud a commentariis 20:25 Sia

were the priests:

20:26 And Ira also the Jairite was a chief ruler about autem scriba Sadoc uero et Abiathar sacerdotes 20:26 Hira

David.

autem Hiaiarites erat sacerdos Dauid

21:1 Then there was a famine in the days of David three 21:1 facta est quoque fames in diebus Dauid tribus annis

years, year after year; and David enquired of the LORD. And the LORD answered, It is for Saul, and for his bloody

house, because he slew the Gibeonites. iugiter et consuluit Dauid oraculum Domini dixitque Do-

minus propter Saul et domum eius et sanguinem quia oc-

21:2 And the king called the Gibeonites, and said unto cidit Gabaonitas 21:2 uocatis ergo Gabaonitis rex dixit ad

them; (now the Gibeonites were not of the children of Israel, but of the remnant of the Amorites; and the children of Israel had sworn unto them: and Saul sought to slay them in his zeal to the children of Israel and Judah.)

eos porro Gabaonitae non sunt de filiis Israhel sed reliquiae Amorreorum filii quippe Israhel iurauerant eis et uoluit Saul

21:3 Wherefore David said unto the Gibeonites, What percutere eos zelo quasi pro filiis Israhel et Iuda 21:3 dixit

shall I do for you? and wherewith shall I make the atonement, that ye may bless the inheritance of the LORD?

ergo Dauid ad Gabaonitas quid faciam uobis et quod erit

21:4 And the Gibeonites said unto him, We will have no uestri piaculum ut benedicatis hereditati Domini 21:4 dixe-

silver nor gold of Saul, nor of his house; neither for us shalt thou kill any man in Israel. And he said, What ye

shall say, that will I do for you. runtque ei Gabaonitae non est nobis super argento et auro quaestio contra Saul et contra domum eius neque uolumus ut interficiatur homo de Israhel ad quos ait quid ergo uultis ut

21:5 And they answered the king, The man that consumed faciam uobis 21:5 qui dixerunt regi uirum qui adtriuit nos et

us, and that devised against us that we should be destroyed from remaining in any of the coasts of Israel,

oppressit inique ita delere debemus ut ne unus quidem resi-

21:6 Let seven men of his sons be delivered unto us, and duus sit de stirpe eius in cunctis finibus Israhel 21:6 dentur

we will hang them up unto the LORD in Gibeah of Saul, whom the LORD did choose. And the king said, I will

give them. nobis septem uiri de filiis eius et crucifigamus eos Domino

in Gabaath Saul quondam electi Domini et ait rex ego dabo

21:7 But the king spared Mephibosheth, the son of 21:7 pepercitque rex Mifiboseth filio Ionathan filii Saul pro-

Jonathan the son of Saul, because of the LORD’s oath that was between them, between David and Jonathan the son

of Saul. pter iusiurandum Domini quod fuerat inter Dauid et inter Io-21:8 But the king took the two sons of Rizpah the daugh- nathan filium Saul 21:8 tulit itaque rex duos filios Respha

ter of Aiah, whom she bare unto Saul, Armoni and Mephibosheth; and the five sons of Michal the daughter of Saul, whom she brought up for Adriel the son of Barzillai the Meholathite:

filiae Ahia quos peperit Saul Armoni et Mifiboseth et quinque filios Michol filiae Saul quos genuerat Hadriheli filio

21:9 And he delivered them into the hands of the Berzellai qui fuit de Molathi 21:9 et dedit eos in manu Ga-

Gibeonites, and they hanged them in the hill before the LORD: and they fell all seven together, and were put to death in the days of harvest, in the first days, in the beginning of barley harvest.

baonitarum qui crucifixerunt illos in monte coram Domino et ceciderunt hii septem simul occisi in diebus messis primis

21:10 And Rizpah the daughter of Aiah took sackcloth, incipiente messione hordei 21:10 tollens autem Respha fi-

and spread it for her upon the rock, from the beginning of harvest until water dropped upon them out of heaven, and suffered neither the birds of the air to rest on them by day, nor the beasts of the field by night.

lia Ahia cilicium substrauit sibi super petram ab initio messis donec stillaret aqua super eos de caelo et non dimisit aues la-

Liber II Samuelis

cerare eos per diem neque bestias per noctem 21:11 et nun- 21:11 And it was told David what Rizpah the daughter of

Aiah, the concubine of Saul, had done.

tiata sunt Dauid quae fecerat Respha filia Ahia concubina Saul 21:12 et abiit Dauid et tulit ossa Saul et ossa Ionathan 21:12 And David went and took the bones of Saul and the

bones of Jonathan his son from the men of Jabeshgilead, which had stolen them from the street of Bethshan, where the Philistines had hanged them, when the Philistines had slain Saul in Gilboa:

filii eius a uiris Iabesgalaad qui furati fuerant ea de platea Bethsan in qua suspenderant eos Philisthim cum interfecissent Saul in Gelboe 21:13 et asportauit inde ossa Saul et 21:13 And he brought up from thence the bones of Saul

and the bones of Jonathan his son; and they gathered the bones of them that were hanged.

ossa Ionathan filii eius et colligentes ossa eorum qui adfixi fuerant 21:14 sepelierunt ea cum ossibus Saul et Ionathan 21:14 And the bones of Saul and Jonathan his son buried

they in the country of Benjamin in Zelah, in the sepulchre of Kish his father: and they performed all that the king commanded. And after that God was intreated for the land.

filii eius in terra Beniamin in latere in sepulchro Cis patris eius feceruntque omnia quae praeceperat rex et repropitiatus

est Deus terrae post haec 21:15 factum est autem rursum 21:15 Moreover the Philistines had yet war again with Israel; and David went down, and his servants with him, and fought against the Philistines: and David waxed faint.

proelium Philisthinorum aduersum Israhel et descendit Dauid et serui eius cum eo et pugnabant contra Philisthim deficiente autem Dauid 21:16 Iesbidenob qui fuit de genere 21:16 And Ishbibenob, which was of the sons of the giant,

the weight of whose spear weighed three hundred shekels of brass in weight, he being girded with a new sword,

Arafa cuius ferrum hastae trecentas uncias adpendebat et thought to have slain David. accinctus erat ense nouo nisus est percutere Dauid 21:17 21:17 But Abishai the son of Zeruiah succoured him, and

smote the Philistine, and killed him. Then the men of David sware unto him, saying, Thou shalt go no more out

praesidioque ei fuit Abisai filius Saruiae et percussum Phi- with us to battle, that thou quench not the light of Israel. listheum interfecit tunc iurauerunt uiri Dauid dicentes non egredieris nobiscum in bellum ne extinguas lucernam Israhel 21:18 secundum quoque fuit bellum in Gob contra Phi- 21:18 And it came to pass after this, that there was again

a battle with the Philistines at Gob: then Sibbechai the Hushathite slew Saph, which was of the sons of the giant.

listheos tunc percussit Sobbochai de Vsathi Seph de stirpe Arafa 21:19 tertium quoque fuit bellum in Gob contra Phi- 21:19 And there was again a battle in Gob with the

Philistines, where Elhanan the son of Jaareoregim, a Bethlehemite, slew the brother of Goliath the Gittite, the staff

listheos in quo percussit Adeodatus filius Saltus polymita- of whose spear was like a weaver’s beam. rius bethleemites Goliath Gettheum cuius hastile hastae erat quasi liciatorium texentium 21:20 quartum bellum fuit in 21:20 And there was yet a battle in Gath, where was a man

of great stature, that had on every hand six fingers, and on every foot six toes, four and twenty in number; and he also

Geth in quo uir excelsus qui senos in manibus pedibusque was born to the giant. habebat digitos id est uiginti et quattuor et erat de origine Arafa 21:21 blasphemauit Israhel percussit autem eum Io- 21:21 And when he defied Israel, Jonathan the son of

Shimeah the brother of David slew him.

nathan filius Sammaa fratris Dauid 21:22 hii quattuor nati 21:22 These four were born to the giant in Gath, and fell

by the hand of David, and by the hand of his servants.

sunt de Arafa in Geth et ceciderunt in manu Dauid et seruorum eius

22:1 locutus est autem Dauid Domino uerba carminis 22:1 And David spake unto the LORD the words of this

song in the day that the LORD had delivered him out of the hand of all his enemies, and out of the hand of Saul:

huius in die qua liberauit eum Dominus de manu omnium

22:2 And he said, The LORD is my rock, and my fortress, inimicorum suorum et de manu Saul 22:2 et ait Dominus and my deliverer;

22:3 The God of my rock; in him will I trust: he is my petra mea et robur meum et saluator meus 22:3 Deus meus

shield, and the horn of my salvation, my high tower, and my refuge, my saviour; thou savest me from violence.

fortis meus sperabo in eum scutum meum et cornu salutis meae eleuator meus et refugium meum saluator meus de in-

22:4 I will call on the LORD, who is worthy to be praised: iquitate liberabis me 22:4 laudabilem inuocabo Dominum

so shall I be saved from mine enemies.

22:5 When the waves of death compassed me, the floods et ab inimicis meis saluus ero 22:5 quia circumdederunt me

of ungodly men made me afraid;

22:6 The sorrows of hell compassed me about; the snares contritiones mortis torrentes Belial terruerunt me 22:6 fu-

of death prevented me;

nes inferi circumdederunt me praeuenerunt me laquei mor-

22:7 In my distress I called upon the LORD, and cried to tis 22:7 in tribulatione mea inuocabo Dominum et ad De-

my God: and he did hear my voice out of his temple, and my cry did enter into his ears.

um meum clamabo et exaudiet de templo suo uocem meam

22:8 Then the earth shook and trembled; the foundations et clamor meus ueniet ad aures eius 22:8 commota est et of heaven moved and shook, because he was wroth.

contremuit terra fundamenta montium concussa sunt et con-

22:9 There went up a smoke out of his nostrils, and fire quassata quoniam iratus est 22:9 ascendit fumus de naribus

out of his mouth devoured: coals were kindled by it.

eius et ignis de ore eius uorauit carbones incensi sunt ab eo

22:10 He bowed the heavens also, and came down; and 22:10 et inclinauit caelos et descendit et caligo sub pedibus

darkness was under his feet.

22:11 And he rode upon a cherub, and did fly: and he was eius 22:11 et ascendit super cherubin et uolauit et lapsus est

seen upon the wings of the wind.

22:12 And he made darkness pavilions round about him, super pinnas uenti 22:12 posuit tenebras in circuitu suo lati-

dark waters, and thick clouds of the skies. 22:13 Through the brightness before him were coals of fire bulum cribrans aquas de nubibus caelorum 22:13 prae fulkindled.

22:14 The LORD thundered from heaven, and the most gore in conspectu eius succensi sunt carbones ignis 22:14

High uttered his voice.

tonabit de caelis Dominus et Excelsus dabit uocem suam

22:15 And he sent out arrows, and scattered them; light- 22:15 misit sagittas et dissipauit eos fulgur et consumpsit eos

ning, and discomfited them.

22:16 And the channels of the sea appeared, the founda- 22:16 et apparuerunt effusiones maris et reuelata sunt fun-

tions of the world were discovered, at the rebuking of the LORD, at the blast of the breath of his nostrils.

damenta orbis ab increpatione Domini ab inspiratione spi-

22:17 He sent from above, he took me; he drew me out of ritus furoris eius 22:17 misit de excelso et adsumpsit me

many waters;

22:18 He delivered me from my strong enemy, and from extraxit me de aquis multis 22:18 liberauit me ab inimico

them that hated me: for they were too strong for me.

meo potentissimo ab his qui oderant me quoniam robustio-

22:19 They prevented me in the day of my calamity: but res me erant 22:19 praeuenit me in die adflictionis meae

the LORD was my stay. 22:20 He brought me forth also into a large place: he de- et factus est Dominus firmamentum meum 22:20 et eduxit livered me, because he delighted in me.

22:21 The LORD rewarded me according to my righteous- me in latitudinem liberauit me quia placuit ei 22:21 retri-

ness: according to the cleanness of my hands hath he recompensed me.

buet mihi Dominus secundum iustitiam meam et secundum

22:22 For I have kept the ways of the LORD, and have not munditiam manuum mearum reddet mihi 22:22 quia custo-

wickedly departed from my God. 22:23 For all his judgments were before me: and as for his diui uias Domini et non egi impie a Deo meo 22:23 omnia statutes, I did not depart from them.

enim iudicia eius in conspectu meo et praecepta eius non

Liber II Samuelis

amoui a me 22:24 et ero perfectus cum eo et custodiam me 22:24 I was also upright before him, and have kept myself

from mine iniquity.

ab iniquitate mea 22:25 et restituet Dominus mihi secun- 22:25 Therefore the LORD hath recompensed me accord-

ing to my righteousness; according to my cleanness in his eye sight.

dum iustitiam meam et secundum munditiam manuum mea-

rum in conspectu oculorum suorum 22:26 cum sancto sanc- 22:26 With the merciful thou wilt shew thyself merciful,

and with the upright man thou wilt shew thyself upright.

tus eris et cum robusto perfectus 22:27 cum electo electus 22:27 With the pure thou wilt shew thyself pure; and with

the froward thou wilt shew thyself unsavoury.

eris et cum peruerso peruerteris 22:28 et populum paupe- 22:28 And the afflicted people thou wilt save: but thine

eyes are upon the haughty, that thou mayest bring them down.

rem saluum facies oculisque tuis excelsos humiliabis 22:29

22:29 For thou art my lamp, O LORD: and the LORD will lighten my darkness.

quia tu lucerna mea Domine et Domine inluminabis tenebras meas 22:30 in te enim curram accinctus in Deo meo trans- 22:30 For by thee I have run through a troop: by my God

have I leaped over a wall.

iliam murum 22:31 Deus inmaculata uia eius eloquium Do- 22:31 As for God, his way is perfect; the word of the

LORD is tried: he is a buckler to all them that trust in him.

mini igne examinatum scutum est omnium sperantium in se 22:32 quis est deus praeter Dominum et quis fortis praeter 22:32 For who is God, save the LORD? and who is a rock,

save our God?

Deum nostrum 22:33 Deus qui accingit me fortitudine et 22:33 God is my strength and power: and he maketh my

way perfect.

conplanauit perfectam uiam meam 22:34 coaequans pedes 22:34 He maketh my feet like hinds’ feet: and setteth me

upon my high places.

meos ceruis et super excelsa mea statuens me 22:35 do- 22:35 He teacheth my hands to war; so that a bow of steel

is broken by mine arms.

cens manus meas ad proelium et conponens quasi arcum aereum brachia mea 22:36 dedisti mihi clypeum salutis tuae 22:36 Thou hast also given me the shield of thy salvation:

and thy gentleness hath made me great.

et mansuetudo mea multiplicauit me 22:37 dilatabis gressus 22:37 Thou hast enlarged my steps under me; so that my

feet did not slip.

meos subtus me et non deficient tali mei 22:38 persequar 22:38 I have pursued mine enemies, and destroyed them;

and turned not again until I had consumed them.

inimicos meos et conteram et non reuertar donec consumam eos 22:39 consumam eos et confringam ut non consurgant 22:39 And I have consumed them, and wounded them, that

they could not arise: yea, they are fallen under my feet.

cadent sub pedibus meis 22:40 accinxisti me fortitudine ad 22:40 For thou hast girded me with strength to battle: them

that rose up against me hast thou subdued under me.

proelium incuruabis resistentes mihi sub me 22:41 inimi- 22:41 Thou hast also given me the necks of mine enemies,

that I might destroy them that hate me.

cos meos dedisti mihi dorsum odientes me et disperdam eos 22:42 clamabunt et non erit qui saluet ad Dominum et 22:42 They looked, but there was none to save; even unto

the LORD, but he answered them not.

non exaudiet eos 22:43 delebo eos ut puluerem terrae quasi 22:43 Then did I beat them as small as the dust of the

earth, I did stamp them as the mire of the street, and did spread them abroad.

lutum platearum comminuam eos atque conpingam 22:44

22:44 Thou also hast delivered me from the strivings of my people, thou hast kept me to be head of the heathen: a people which I knew not shall serve me.

saluabis me a contradictionibus populi mei custodies in ca-

put gentium populus quem ignoro seruiet mihi 22:45 filii 22:45 Strangers shall submit themselves unto me: as soon

as they hear, they shall be obedient unto me.

alieni resistent mihi auditu auris oboedient mihi 22:46 fi- 22:46 Strangers shall fade away, and they shall be afraid

out of their close places.

lii alieni defluxerunt et contrahentur in angustiis suis 22:47 22:47 The LORD liveth; and blessed be my rock; and ex-

alted be the God of the rock of my salvation.

uiuit Dominus et benedictus Deus meus et exaltabitur Deus fortis salutis meae 22:48 Deus qui das uindictas mihi et de- 22:48 It is God that avengeth me, and that bringeth down

the people under me.

Biblia Sacra Vulgata

22:49 And that bringeth me forth from mine enemies: thou icis populos sub me 22:49 qui educis me ab inimicis meis also hast lifted me up on high above them that rose up against me: thou hast delivered me from the violent man.

et a resistentibus mihi eleuas me a uiro iniquo liberabis me

22:50 Therefore I will give thanks unto thee, O LORD, 22:50 propterea confitebor tibi Domine in gentibus et nomini

among the heathen, and I will sing praises unto thy name.

22:51 He is the tower of salvation for his king: and tuo cantabo 22:51 magnificanti salutes regis sui et facienti

sheweth mercy to his anointed, unto David, and to his seed for evermore.

misericordiam christo suo Dauid et semini eius in sempiternum

23:1 Now these be the last words of David. David the son 23:1 haec autem sunt uerba nouissima quae dixit Dauid of Jesse said, and the man who was raised up on high, the anointed of the God of Jacob, and the sweet psalmist of

Israel, said, filius Isai dixit uir cui constitutum est de christo Dei Iacob 23:2 The Spirit of the LORD spake by me, and his word egregius psalta Israhel 23:2 spiritus Domini locutus est per

was in my tongue. 23:3 The God of Israel said, the Rock of Israel spake to me et sermo eius per linguam meam 23:3 dixit Deus Israme, He that ruleth over men must be just, ruling in the fear of God.

hel mihi locutus est Fortis Israhel dominator hominum iustus

23:4 And he shall be as the light of the morning, when the dominator in timore Dei 23:4 sicut lux aurorae oriente sole

sun riseth, even a morning without clouds; as the tender grass springing out of the earth by clear shining after rain.

mane absque nubibus rutilat et sicut pluuiis germinat herba

23:5 Although my house be not so with God; yet he de terra 23:5 nec tanta est domus mea apud Deum ut pac-

hath made with me an everlasting covenant, ordered in all things, and sure: for this is all my salvation, and all my

desire, although he make it not to grow. tum aeternum iniret mecum firmum in omnibus atque muni-

tum cuncta enim salus mea et omnis uoluntas nec est quic-

23:6 But the sons of Belial shall be all of them as thorns quam ex ea quod non germinet 23:6 praeuaricatores autem

thrust away, because they cannot be taken with hands:

quasi spinae euellentur uniuersi quae non tolluntur manibus

23:7 But the man that shall touch them must be fenced 23:7 et si quis tangere uoluerit eas armabitur ferro et ligno

with iron and the staff of a spear; and they shall be utterly burned with fire in the same place.

lanceato igneque succensae conburentur usque ad nihilum

23:8 These be the names of the mighty men whom David 23:8 haec nomina fortium Dauid Sedens in cathedra sapien-

had: The Tachmonite that sat in the seat, chief among the captains; the same was Adino the Eznite: he lift up his

spear against eight hundred, whom he slew at one time. tissimus princeps inter tres ipse est quasi tenerrimus ligni 23:9 And after him was Eleazar the son of Dodo the Aho- uermiculus qui octingentos interfecit impetu uno 23:9 post

hite, one of the three mighty men with David, when they defied the Philistines that were there gathered together to

battle, and the men of Israel were gone away: hunc Eleazar filius patrui eius Ahoi inter tres fortes qui erant

cum Dauid quando exprobrauerunt Philisthim et congregati

23:10 He arose, and smote the Philistines until his hand sunt illuc in proelium 23:10 cumque ascendissent uiri Isra-

was weary, and his hand clave unto the sword: and the LORD wrought a great victory that day; and the people

returned after him only to spoil. hel ipse stetit et percussit Philistheos donec deficeret manus eius et obrigesceret cum gladio fecitque Dominus salutem magnam in die illa et populus qui fugerat reuersus est ad cae-

23:11 And after him was Shammah the son of Agee the sorum spolia detrahenda 23:11 et post hunc Semma filius

Hararite. And the Philistines were gathered together into a troop, where was a piece of ground full of lentiles: and

the people fled from the Philistines. Age de Arari et congregati sunt Philisthim in statione erat

quippe ibi ager plenus lente cumque fugisset populus a facie

23:12 But he stood in the midst of the ground, and de- Philisthim 23:12 stetit ille in medio agri et tuitus est eum

fended it, and slew the Philistines: and the LORD wrought a great victory.

Liber II Samuelis

percussitque Philistheos et fecit Dominus salutem magnam 23:13 necnon ante descenderant tres qui erant principes inter 23:13 And three of the thirty chief went down, and came

to David in the harvest time unto the cave of Adullam: and the troop of the Philistines pitched in the valley of

triginta et uenerant tempore messis ad Dauid in speluncam Rephaim. Odollam castra autem Philisthim erant posita in ualle Gigantum 23:14 et Dauid erat in praesidio porro statio Philisthi- 23:14 And David was then in an hold, and the garrison of

the Philistines was then in Bethlehem.

norum tunc erat in Bethleem 23:15 desiderauit igitur Dauid 23:15 And David longed, and said, Oh that one would give

me drink of the water of the well of Bethlehem, which is by the gate!

et ait si quis mihi daret potum aquae de cisterna quae est in Bethleem iuxta portam 23:16 inruperunt ergo tres fortes 23:16 And the three mighty men brake through the host

of the Philistines, and drew water out of the well of Bethlehem, that was by the gate, and took it, and brought it to David: nevertheless he would not drink thereof, but poured it out unto the LORD.

castra Philisthinorum et hauserunt aquam de cisterna Bethleem quae erat iuxta portam et adtulerunt ad Dauid at ille noluit bibere sed libauit illam Domino 23:17 dicens propitius 23:17 And he said, Be it far from me, O LORD, that I

should do this: is not this the blood of the men that went in jeopardy of their lives? therefore he would not drink it.

mihi sit Dominus ne faciam hoc num sanguinem hominum These things did these three mighty men. istorum qui profecti sunt et animarum periculum bibam noluit ergo bibere haec fecerunt tres robustissimi 23:18 Abi- 23:18 And Abishai, the brother of Joab, the son of Zeruiah,

was chief among three. And he lifted up his spear against three hundred, and slew them, and had the name among

sai quoque frater Ioab filius Saruiae princeps erat de tribus three. ipse est qui eleuauit hastam suam contra trecentos quos interfecit nominatus in tribus 23:19 et inter tres nobilior erat- 23:19 Was he not most honourable of three? therefore he

was their captain: howbeit he attained not unto the first three.

que eorum princeps sed usque ad tres primos non peruenerat 23:20 et Banaias filius Ioiada uiri fortissimi magnorum ope- 23:20 And Benaiah the son of Jehoiada, the son of a

valiant man, of Kabzeel, who had done many acts, he slew two lionlike men of Moab: he went down also and slew a

rum de Capsehel ipse percussit duos leones Moab et ipse lion in the midst of a pit in time of snow: descendit et percussit leonem in media cisterna diebus niuis 23:21 ipse quoque interfecit uirum aegyptium uirum dignum 23:21 And he slew an Egyptian, a goodly man: and the

Egyptian had a spear in his hand; but he went down to him with a staff, and plucked the spear out of the Egyptian’s

spectaculo habentem in manu hastam itaque cum descendis- hand, and slew him with his own spear. set ad eum in uirga ui extorsit hastam de manu Aegyptii et interfecit eum hasta sua 23:22 haec fecit Banaias filius Ioia- 23:22 These things did Benaiah the son of Jehoiada, and

had the name among three mighty men.

dae 23:23 et ipse nominatus inter tres robustos qui erant in- 23:23 He was more honourable than the thirty, but he at-

tained not to the first three. And David set him over his guard.

ter triginta nobiliores uerumtamen usque ad tres non peruenerat fecitque eum Dauid sibi auricularium a secreto 23:24 23:24 Asahel the brother of Joab was one of the thirty;

Elhanan the son of Dodo of Bethlehem,

Asahel frater Ioab inter triginta Eleanan filius patrui eius de Bethleem 23:25 Semma de Arari Helica de Arodi 23:26 23:25 Shammah the Harodite, Elika the Harodite,

23:26 Helez the Paltite, Ira the son of Ikkesh the Tekoite,

Helas de Felthi Hira filius Aces de Thecua 23:27 Abiezer de 23:27 Abiezer the Anethothite, Mebunnai the Hushathite,

Anathoth Mobonnai de Vsathi 23:28 Selmon Aohites Ma- 23:28 Zalmon the Ahohite, Maharai the Netophathite, harai Netophathites 23:29 Heled filius Banaa et ipse Neto- 23:29 Heleb the son of Baanah, a Netophathite, Ittai the

son of Ribai out of Gibeah of the children of Benjamin,

Biblia Sacra Vulgata

phathites Hithai filius Ribai de Gebeeth filiorum Beniamin

23:30 Benaiah the Pirathonite, Hiddai of the brooks of 23:30 Banahi Aufrathonites Heddai de torrente Gaas 23:31

Gaash, 23:31 Abialbon the Arbathite, Azmaveth the Barhumite,

Abialbon Arbathites Azmaueth de Beromi 23:32 Eliaba de

23:32 Eliahba the Shaalbonite, of the sons of Jashen, Jonathan,

Salboni filii Iasen Ionathan 23:33 Semma de Horodi Haiam

23:33 Shammah the Hararite, Ahiam the son of Sharar the Hararite,

23:34 Eliphelet the son of Ahasbai, the son of the filius Sarar Arorites 23:34 Elifeleth filius Aasbai filii Maa-

Maachathite, Eliam the son of Ahithophel the Gilonite,

23:35 Hezrai the Carmelite, Paarai the Arbite, chathi Heliam filius Ahitofel Gelonites 23:35 Esrai de Car-23:36 Igal the son of Nathan of Zobah, Bani the Gadite, melo Farai de Arbi 23:36 Igaal filius Nathan de Soba Bonni 23:37 Zelek the Ammonite, Nahari the Beerothite, ar- de Gaddi 23:37 Selech de Ammoni Naharai Berothites ar-

mourbearer to Joab the son of Zeruiah,

23:38 Ira an Ithrite, Gareb an Ithrite, miger Ioab filii Saruiae 23:38 Hira Hiethrites Gareb et ipse 23:39 Uriah the Hittite: thirty and seven in all. Hiethrites 23:39 Vrias Hettheus omnes triginta septem

24:1 And again the anger of the LORD was kindled 24:1 et addidit furor Domini irasci contra Israhel commo-

against Israel, and he moved David against them to say, Go, number Israel and Judah.

uitque Dauid in eis dicentem uade numera Israhel et Iudam

24:2 For the king said to Joab the captain of the host, 24:2 dixitque rex ad Ioab principem exercitus sui perambula

which was with him, Go now through all the tribes of Israel, from Dan even to Beersheba, and number ye the

people, that I may know the number of the people. omnes tribus Israhel a Dan usque Bersabee et numerate po-24:3 And Joab said unto the king, Now the LORD thy pulum ut sciam numerum eius 24:3 dixitque Ioab regi ad-

God add unto the people, how many soever they be, an hundredfold, and that the eyes of my lord the king may see it: but why doth my lord the king delight in this thing? augeat Dominus Deus tuus ad populum quantus nunc est ite-

rumque centuplicet in conspectu domini mei regis sed quid

24:4 Notwithstanding the king’s word prevailed against sibi dominus meus rex uult in re huiuscemodi 24:4 obtinuit

Joab, and against the captains of the host. And Joab and the captains of the host went out from the presence of the

king, to number the people of Israel. autem sermo regis uerba Ioab et principum exercitus egres-

susque est Ioab et principes militum a facie regis ut nume-

24:5 And they passed over Jordan, and pitched in Aroer, rarent populum Israhel 24:5 cumque pertransissent Iorda-

on the right side of the city that lieth in the midst of the river of Gad, and toward Jazer:

nem uenerunt in Aroer ad dextram urbis quae est in ualle

24:6 Then they came to Gilead, and to the land of Tah- Gad 24:6 et per Iazer transierunt in Galaad et in terram in-

timhodshi; and they came to Danjaan, and about to Zidon,

feriorem Hodsi et uenerunt in Dan siluestria circumeuntes-

24:7 And came to the strong hold of Tyre, and to all the que iuxta Sidonem 24:7 transierunt propter moenia Tyri et

cities of the Hivites, and of the Canaanites: and they went out to the south of Judah, even to Beersheba.

omnem terram Heuei et Chananei ueneruntque ad meridiem

24:8 So when they had gone through all the land, they Iuda in Bersabee 24:8 et lustrata uniuersa terra adfuerunt

came to Jerusalem at the end of nine months and twenty days.

post nouem menses et uiginti dies in Hierusalem 24:9 de-

24:9 And Joab gave up the sum of the number of the people unto the king: and there were in Israel eight hundred thousand valiant men that drew the sword; and the men of Judah were five hundred thousand men.

dit ergo Ioab numerum descriptionis populi regi et inuenta sunt de Israhel octingenta milia uirorum fortium qui educe-

24:10 And David’s heart smote him after that he had num- rent gladium et de Iuda quingenta milia pugnatorum 24:10

bered the people. And David said unto the LORD, I have sinned greatly in that I have done: and now, I beseech thee, O LORD, take away the iniquity of thy servant; for I have done very foolishly.

percussit autem cor Dauid eum postquam numeratus est populus et dixit Dauid ad Dominum peccaui ualde in hoc facto

Liber II Samuelis

sed precor Domine ut transferas iniquitatem serui tui quia stulte egi nimis 24:11 surrexit itaque Dauid mane et sermo 24:11 For when David was up in the morning, the word

of the LORD came unto the prophet Gad, David’s seer, saying,

Domini factus est ad Gad propheten et uidentem Dauid dicens 24:12 uade et loquere ad Dauid haec dicit Dominus 24:12 Go and say unto David, Thus saith the LORD, I

offer thee three things; choose thee one of them, that I may do it unto thee.

trium tibi datur optio elige unum quod uolueris ex his ut faciam tibi 24:13 cumque uenisset Gad ad Dauid nuntiauit ei 24:13 So Gad came to David, and told him, and said unto

him, Shall seven years of famine come unto thee in thy land? or wilt thou flee three months before thine enemies, while they pursue thee? or that there be three days’ pestilence in thy land? now advise, and see what answer I shall return to him that sent me.

dicens aut septem annis ueniet tibi fames in terra tua aut tribus mensibus fugies aduersarios tuos et illi persequentur aut certe tribus diebus erit pestilentia in terra tua nunc ergo delibera et uide quem respondeam ei qui me misit sermonem 24:14 dixit autem Dauid ad Gad artor nimis sed melius est 24:14 And David said unto Gad, I am in a great strait: let

us fall now into the hand of the LORD; for his mercies are great: and let me not fall into the hand of man.

ut incidam in manu Domini multae enim misericordiae eius sunt quam in manu hominis 24:15 inmisitque Dominus pes- 24:15 So the LORD sent a pestilence upon Israel from

the morning even to the time appointed: and there died of the people from Dan even to Beersheba seventy thousand

tilentiam in Israhel de mane usque ad tempus constitutum et men. mortui sunt ex populo a Dan usque Bersabee septuaginta milia uirorum 24:16 cumque extendisset manum angelus Dei 24:16 And when the angel stretched out his hand upon

Jerusalem to destroy it, the LORD repented him of the evil, and said to the angel that destroyed the people, It is enough: stay now thine hand. And the angel of the LORD was by the threshingplace of Araunah the Jebusite.

super Hierusalem ut disperderet eam misertus est Dominus super adflictione et ait angelo percutienti populum sufficit nunc contine manum tuam erat autem angelus Domini iuxta aream Areuna Iebusei 24:17 dixitque Dauid ad Dominum 24:17 And David spake unto the LORD when he saw the

angel that smote the people, and said, Lo, I have sinned, and I have done wickedly: but these sheep, what have they done? let thine hand, I pray thee, be against me, and against my father’s house.

cum uidisset angelum caedentem populum ego sum qui peccaui ego inique egi isti qui oues sunt quid fecerunt uertatur obsecro manus tua contra me et contra domum patris mei 24:18 uenit autem Gad ad Dauid in die illa et dixit ei ascende 24:18 And Gad came that day to David, and said unto him,

Go up, rear an altar unto the LORD in the threshingfloor of Araunah the Jebusite.

constitue Domino altare in area Areuna Iebusei 24:19 et as-

24:19 And David, according to the saying of Gad, went up as the LORD commanded.

cendit Dauid iuxta sermonem Gad quem praeceperat ei Dominus 24:20 conspiciensque Areuna animaduertit regem et 24:20 And Araunah looked, and saw the king and his

servants coming on toward him: and Araunah went out, and bowed himself before the king on his face upon the

seruos eius transire ad se 24:21 et egressus adorauit regem ground.

24:21 And Araunah said, Wherefore is my lord the king come to his servant? And David said, To buy the threshingfloor of thee, to build an altar unto the LORD, that the plague may be stayed from the people.

prono uultu in terra et ait quid causae est ut ueniat dominus meus rex ad seruum suum cui Dauid ait ut emam a te aream et aedificem altare Domino et cesset interfectio quae grassatur in populo 24:22 et ait Areuna ad Dauid accipiat 24:22 And Araunah said unto David, Let my lord the king

take and offer up what seemeth good unto him: behold, here be oxen for burnt sacrifice, and threshing instruments

et offerat dominus meus rex sicut ei placet habes boues in and other instruments of the oxen for wood.

Biblia Sacra Vulgata

holocaustum et plaustrum et iuga boum in usum lignorum

24:23 All these things did Araunah, as a king, give unto 24:23 omnia dedit Areuna rex regi dixitque Areuna ad re-

the king. And Araunah said unto the king, The LORD thy God accept thee.

gem Dominus Deus tuus suscipiat uotum tuum 24:24 cui

24:24 And the king said unto Araunah, Nay; but I will surely buy it of thee at a price: neither will I offer burnt offerings unto the LORD my God of that which doth cost me nothing. So David bought the threshingfloor and the oxen for fifty shekels of silver.

respondens rex ait nequaquam ut uis sed emam pretio a te et non offeram Domino Deo meo holocausta gratuita emit ergo

24:25 And David built there an altar unto the LORD, and Dauid aream et boues argenti siclis quinquaginta 24:25 et

offered burnt offerings and peace offerings. So the LORD was intreated for the land, and the plague was stayed from

\underline{Liber I Regum}

Israel. aedificauit ibi Dauid altare Domino et obtulit holocausta et pacifica et repropitiatus est Dominus terrae et cohibita est plaga ab Israhel 1:1 et rex Dauid senuerat habebatque aetatis plurimos 1:1 Now king David was old and stricken in years; and

they covered him with clothes, but he gat no heat.

dies cumque operiretur uestibus non calefiebat 1:2 dixerunt 1:2 Wherefore his servants said unto him, Let there be

sought for my lord the king a young virgin: and let her stand before the king, and let her cherish him, and let her

ergo ei serui sui quaeramus domino nostro regi adulescentu- lie in thy bosom, that my lord the king may get heat. lam uirginem et stet coram rege et foueat eum dormiatque in sinu tuo et calefaciat dominum nostrum regem 1:3 quae- 1:3 So they sought for a fair damsel throughout all the

coasts of Israel, and found Abishag a Shunammite, and brought her to the king.

sierunt igitur adulescentulam speciosam in omnibus finibus Israhel et inuenerunt Abisag Sunamitin et adduxerunt eam ad regem 1:4 erat autem puella pulchra nimis dormiebat- 1:4 And the damsel was very fair, and cherished the king,

and ministered to him: but the king knew her not.

que cum rege et ministrabat ei rex uero non cognouit eam 1:5 Adonias autem filius Aggith eleuabatur dicens ego reg- 1:5 Then Adonijah the son of Haggith exalted himself,

saying, I will be king: and he prepared him chariots and horsemen, and fifty men to run before him.

nabo fecitque sibi currum et equites et quinquaginta uiros qui ante eum currerent 1:6 nec corripuit eum pater suus 1:6 And his father had not displeased him at any time in

saying, Why hast thou done so? and he also was a very goodly man; and his mother bare him after Absalom.

aliquando dicens quare hoc fecisti erat autem et ipse pulcher ualde secundus natu post Absalom 1:7 et sermo ei cum 1:7 And he conferred with Joab the son of Zeruiah, and

with Abiathar the priest: and they following Adonijah helped him.

Ioab filio Saruiae et cum Abiathar sacerdote qui adiuuabant partes Adoniae 1:8 Sadoc uero sacerdos et Banaias filius 1:8 But Zadok the priest, and Benaiah the son of Je-

hoiada, and Nathan the prophet, and Shimei, and Rei, and the mighty men which belonged to David, were not with

Ioiadae et Nathan propheta et Semei et Rhei et robur exer- Adonijah. citus Dauid non erat cum Adonia 1:9 immolatis ergo Ad- 1:9 And Adonijah slew sheep and oxen and fat cattle by

the stone of Zoheleth, which is by Enrogel, and called all his brethren the king’s sons, and all the men of Judah the

onias arietibus et uitulis et uniuersis pinguibus iuxta lapidem king’s servants: Zoheleth qui erat uicinus fonti Rogel uocauit uniuersos fratres suos filios regis et omnes uiros Iuda seruos regis 1:10 1:10 But Nathan the prophet, and Benaiah, and the mighty

men, and Solomon his brother, he called not.

Nathan autem prophetam et Banaiam et robustos quosque et Salomonem fratrem suum non uocauit 1:11 dixit itaque 1:11 Wherefore Nathan spake unto Bathsheba the mother

of Solomon, saying, Hast thou not heard that Adonijah the son of Haggith doth reign, and David our lord knoweth it

Nathan ad Bethsabee matrem Salomonis num audisti quod not? regnauerit Adonias filius Aggith et dominus noster Dauid hoc ignorat 1:12 nunc ergo ueni accipe a me consilium et 1:12 Now therefore come, let me, I pray thee, give thee

counsel, that thou mayest save thine own life, and the life of thy son Solomon.

salua animam tuam filiique tui Salomonis 1:13 uade et in-

1:13 Go and get thee in unto king David, and say unto him, Didst not thou, my lord, O king, swear unto thine handmaid, saying, Assuredly Solomon thy son shall reign after me, and he shall sit upon my throne? why then doth Adonijah reign?

gredere ad regem Dauid et dic ei nonne tu domine mi rex iurasti mihi ancillae tuae dicens quod Salomon filius tuus regnabit post me et ipse sedebit in solio meo quare ergo re-

gnauit Adonias 1:14 et adhuc ibi te loquente cum rege ego 1:14 Behold, while thou yet talkest there with the king, I

also will come in after thee, and confirm thy words.

ueniam post te et conplebo sermones tuos 1:15 ingressa est 1:15 And Bathsheba went in unto the king into the cham-

ber: and the king was very old; and Abishag the Shunammite ministered unto the king.

itaque Bethsabee ad regem in cubiculo rex autem senuerat

Biblia Sacra Vulgata

1:16 And Bathsheba bowed, and did obeisance unto the nimis et Abisag Sunamitis ministrabat ei 1:16 inclinauit se

king. And the king said, What wouldest thou?

Bethsabee et adorauit regem ad quam rex quid tibi inquit

1:17 And she said unto him, My lord, thou swarest by the uis 1:17 quae respondens ait domine mi tu iurasti per Do-

LORD thy God unto thine handmaid, saying, Assuredly Solomon thy son shall reign after me, and he shall sit upon

my throne. minum Deum tuum ancillae tuae Salomon filius tuus regna-1:18 And now, behold, Adonijah reigneth; and now, my bit post me et ipse sedebit in solio meo 1:18 et ecce nunc

lord the king, thou knowest it not:

1:19 And he hath slain oxen and fat cattle and sheep in Adonias regnauit te domine mi rex ignorante 1:19 macta-

abundance, and hath called all the sons of the king, and Abiathar the priest, and Joab the captain of the host: but

Solomon thy servant hath he not called. uit boues et pinguia quaeque et arietes plurimos et uocauit omnes filios regis Abiathar quoque sacerdotem et Ioab principem militiae Salomonem autem seruum tuum non uoca-

1:20 And thou, my lord, O king, the eyes of all Israel are uit 1:20 uerumtamen domine mi rex in te oculi respiciunt

upon thee, that thou shouldest tell them who shall sit on the throne of my lord the king after him.

totius Israhel ut indices eis qui sedere debeat in solio tuo

1:21 Otherwise it shall come to pass, when my lord domine mi rex post te 1:21 eritque cum dormierit dominus

the king shall sleep with his fathers, that I and my son Solomon shall be counted offenders.

meus rex cum patribus suis erimus ego et filius meus Salo-

1:22 And, lo, while she yet talked with the king, Nathan mon peccatores 1:22 adhuc illa loquente cum rege Nathan

the prophet also came in.

1:23 And they told the king, saying, Behold Nathan the prophetes uenit 1:23 et nuntiauerunt regi dicentes adest Na-

prophet. And when he was come in before the king, he bowed himself before the king with his face to the ground.

than propheta cumque introisset ante conspectum regis et

1:24 And Nathan said, My lord, O king, hast thou said, adorasset eum pronus in terram 1:24 dixit Nathan domine

Adonijah shall reign after me, and he shall sit upon my throne?

mi rex tu dixisti Adonias regnet post me et ipse sedeat super

1:25 For he is gone down this day, and hath slain oxen and thronum meum 1:25 quia descendit hodie et immolauit bo-

fat cattle and sheep in abundance, and hath called all the king’s sons, and the captains of the host, and Abiathar the priest; and, behold, they eat and drink before him, and say, God save king Adonijah.

ues et pinguia et arietes plurimos et uocauit uniuersos filios regis et principes exercitus Abiathar quoque sacerdotem illisque uescentibus et bibentibus coram eo et dicentibus uiuat

1:26 But me, even me thy servant, and Zadok the priest, rex Adonias 1:26 me seruum tuum et Sadoc sacerdotem et and Benaiah the son of Jehoiada, and thy servant Solomon, hath he not called.

Banaiam filium Ioiadae et Salomonem famulum tuum non

1:27 Is this thing done by my lord the king, and thou uocauit 1:27 numquid a domino meo rege exiuit hoc uer-

hast not shewed it unto thy servant, who should sit on the throne of my lord the king after him?

bum et mihi non indicasti seruo tuo qui sessurus esset super

1:28 Then king David answered and said, Call me thronum domini mei regis post eum 1:28 et respondit rex

Bathsheba. And she came into the king’s presence, and stood before the king.

Dauid dicens uocate ad me Bethsabee quae cum fuisset in-

1:29 And the king sware, and said, As the LORD liveth, gressa coram rege et stetisset ante eum 1:29 iurauit rex et

that hath redeemed my soul out of all distress,

ait uiuit Dominus qui eruit animam meam de omni angustia

1:30 Even as I sware unto thee by the LORD God of Israel, 1:30 quia sicut iuraui tibi per Dominum Deum Israhel dicens

saying, Assuredly Solomon thy son shall reign after me, and he shall sit upon my throne in my stead; even so will

I certainly do this day. Salomon filius tuus regnabit post me et ipse sedebit super

1:31 Then Bathsheba bowed with her face to the earth, and solium meum pro me sic faciam hodie 1:31 submissoque did reverence to the king, and said, Let my lord king David live for ever.

Bethsabee in terram uultu adorauit regem dicens uiuat dominus meus rex Dauid in aeternum 1:32 dixit quoque rex 1:32 And king David said, Call me Zadok the priest, and

Nathan the prophet, and Benaiah the son of Jehoiada. And they came before the king.

Dauid uocate mihi Sadoc sacerdotem et Nathan propheten et Banaiam filium Ioiadae qui cum ingressi fuissent coram rege 1:33 dixit ad eos tollite uobiscum seruos domini uestri 1:33 The king also said unto them, Take with you the ser-

vants of your lord, and cause Solomon my son to ride upon mine own mule, and bring him down to Gihon:

et inponite Salomonem filium meum super mulam meam et ducite eum in Gion 1:34 et unguat eum ibi Sadoc sacer- 1:34 And let Zadok the priest and Nathan the prophet

anoint him there king over Israel: and blow ye with the trumpet, and say, God save king Solomon.

dos et Nathan propheta in regem super Israhel et canetis bucina atque dicetis uiuat rex Salomon 1:35 et ascendetis 1:35 Then ye shall come up after him, that he may come

and sit upon my throne; for he shall be king in my stead: and I have appointed him to be ruler over Israel and over

post eum et ueniet et sedebit super solium meum et ipse re- Judah. gnabit pro me illique praecipiam ut sit dux super Israhel et super Iudam 1:36 et respondit Banaias filius Ioiadae regi 1:36 And Benaiah the son of Jehoiada answered the king,

and said, Amen: the LORD God of my lord the king say so too.

dicens amen sic loquatur Dominus Deus domini mei regis 1:37 quomodo fuit Dominus cum domino meo rege sic sit 1:37 As the LORD hath been with my lord the king, even

so be he with Solomon, and make his throne greater than the throne of my lord king David.

cum Salomone et sublimius faciat solium eius a solio domini mei regis Dauid 1:38 descendit ergo Sadoc sacerdos 1:38 So Zadok the priest, and Nathan the prophet, and Be-

naiah the son of Jehoiada, and the Cherethites, and the Pelethites, went down, and caused Solomon to ride upon

et Nathan propheta et Banaias filius Ioiadae et Cherethi et king David’s mule, and brought him to Gihon. Felethi et inposuerunt Salomonem super mulam regis Dauid et adduxerunt eum in Gion 1:39 sumpsitque Sadoc sacer- 1:39 And Zadok the priest took an horn of oil out of the

tabernacle, and anointed Solomon. And they blew the trumpet; and all the people said, God save king Solomon.

dos cornu olei de tabernaculo et unxit Salomonem et cecinerunt bucina et dixit omnis populus uiuat rex Salomon 1:40 1:40 And all the people came up after him, and the people

piped with pipes, and rejoiced with great joy, so that the earth rent with the sound of them.

et ascendit uniuersa multitudo post eum et populus canentium tibiis et laetantium gaudio magno et insonuit terra ad clamorem eorum 1:41 audiuit autem Adonias et omnes qui 1:41 And Adonijah and all the guests that were with him

heard it as they had made an end of eating. And when Joab heard the sound of the trumpet, he said, Wherefore is this

inuitati fuerant ab eo iamque conuiuium finitum erat sed et noise of the city being in an uproar? Ioab audita uoce tubae ait quid sibi uult clamor ciuitatis tumultuantis 1:42 adhuc illo loquente Ionathan filius Abiat- 1:42 And while he yet spake, behold, Jonathan the son

of Abiathar the priest came; and Adonijah said unto him, Come in; for thou art a valiant man, and bringest good

har sacerdotis uenit cui dixit Adonias ingredere quia uir for- tidings. tis es et bona nuntians 1:43 responditque Ionathan Adoniae 1:43 And Jonathan answered and said to Adonijah, Verily

our lord king David hath made Solomon king.

nequaquam dominus enim noster rex Dauid regem constituit Salomonem 1:44 misitque cum eo Sadoc sacerdotem et 1:44 And the king hath sent with him Zadok the priest, and

Nathan the prophet, and Benaiah the son of Jehoiada, and the Cherethites, and the Pelethites, and they have caused

Nathan prophetam et Banaiam filium Ioiadae et Cherethi et him to ride upon the king’s mule: Felethi et inposuerunt eum super mulam regis 1:45 unxe- 1:45 And Zadok the priest and Nathan the prophet have

anointed him king in Gihon: and they are come up from thence rejoicing, so that the city rang again. This is the noise that ye have heard.

Biblia Sacra Vulgata

runtque eum Sadoc sacerdos et Nathan propheta regem in Gion et ascenderunt inde laetantes et insonuit ciuitas haec

1:46 And also Solomon sitteth on the throne of the king- est uox quam audistis 1:46 sed et Salomon sedit super so-

dom.

1:47 And moreover the king’s servants came to bless our lio regni 1:47 et ingressi serui regis benedixerunt domino

lord king David, saying, God make the name of Solomon better than thy name, and make his throne greater than thy

throne. And the king bowed himself upon the bed. nostro regi Dauid dicentes amplificet Deus nomen Salomo-

nis super nomen tuum et magnificet thronum eius super th-

1:48 And also thus said the king, Blessed be the LORD ronum tuum et adorauit rex in lectulo suo 1:48 insuper et

God of Israel, which hath given one to sit on my throne this day, mine eyes even seeing it.

haec locutus est benedictus Dominus Deus Israhel qui dedit

1:49 And all the guests that were with Adonijah were hodie sedentem in solio meo uidentibus oculis meis 1:49

afraid, and rose up, and went every man his way.

territi sunt ergo et surrexerunt omnes qui inuitati fuerant ab

1:50 And Adonijah feared because of Solomon, and arose, Adonia et iuit unusquisque in uiam suam 1:50 Adonias auand went, and caught hold on the horns of the altar.

tem timens Salomonem surrexit et abiit tenuitque cornu al-

1:51 And it was told Solomon, saying, Behold, Adonijah taris 1:51 et nuntiauerunt Salomoni dicentes ecce Adonias

feareth king Solomon: for, lo, he hath caught hold on the horns of the altar, saying, Let king Solomon swear unto

me today that he will not slay his servant with the sword. timens regem Salomonem tenuit cornu altaris dicens iuret

mihi hodie rex Salomon quod non interficiat seruum suum

1:52 And Solomon said, If he will shew himself a worthy gladio 1:52 dixitque Salomon si fuerit uir bonus non cadet

man, there shall not an hair of him fall to the earth: but if wickedness shall be found in him, he shall die.

ne unus quidem capillus eius in terram sin autem malum in-

1:53 So king Solomon sent, and they brought him down uentum fuerit in eo morietur 1:53 misit ergo rex Salomon et

from the altar. And he came and bowed himself to king Solomon: and Solomon said unto him, Go to thine house.

eduxit eum ab altari et ingressus adorauit regem Salomonem dixitque ei Salomon uade in domum tuam

2:1 Now the days of David drew nigh that he should die; 2:1 adpropinquauerant autem dies Dauid ut moreretur

and he charged Solomon his son, saying, 2:2 I go the way of all the earth: be thou strong therefore, praecepitque Salomoni filio suo dicens 2:2 ego ingredior and shew thyself a man;

2:3 And keep the charge of the LORD thy God, to walk uiam uniuersae terrae confortare et esto uir 2:3 et obserua

in his ways, to keep his statutes, and his commandments, and his judgments, and his testimonies, as it is written in the law of Moses, that thou mayest prosper in all that thou doest, and whithersoever thou turnest thyself:

custodias Domini Dei tui ut ambules in uiis eius et custodias caerimonias eius et praecepta eius et iudicia et testimonia sicut scriptum est in lege Mosi ut intellegas uniuersa quae

2:4 That the LORD may continue his word which he spake facis et quocumque te uerteris 2:4 ut confirmet Dominus

concerning me, saying, If thy children take heed to their way, to walk before me in truth with all their heart and with all their soul, there shall not fail thee (said he) a man on the throne of Israel.

sermones suos quos locutus est de me dicens si custodierint filii tui uiam suam et ambulauerint coram me in ueritate in omni corde suo et in omni anima sua non auferetur tibi uir de

2:5 Moreover thou knowest also what Joab the son of solio Israhel 2:5 tu quoque nosti quae fecerit mihi Ioab fi-

Zeruiah did to me, and what he did to the two captains of the hosts of Israel, unto Abner the son of Ner, and unto Amasa the son of Jether, whom he slew, and shed the blood of war in peace, and put the blood of war upon his girdle that was about his loins, and in his shoes that were on his feet.

lius Saruiae quae fecerit duobus principibus exercitus Israhel Abner filio Ner et Amasa filio Iether quos occidit et effudit sanguinem belli in pace et posuit cruorem proelii in balteo suo qui erat circa lumbos eius et in calciamento suo quod erat in pedibus eius 2:6 facies ergo iuxta sapientiam tuam 2:6 Do therefore according to thy wisdom, and let not his

hoar head go down to the grave in peace.

et non deduces canitiem eius pacifice ad inferos 2:7 sed et 2:7 But shew kindness unto the sons of Barzillai the Gileadite, and let them be of those that eat at thy table: for so they came to me when I fled because of Absalom thy

filiis Berzellai Galaaditis reddes gratiam eruntque comeden- brother. tes in mensa tua occurrerunt enim mihi quando fugiebam a facie Absalom fratris tui 2:8 habes quoque apud te Semei 2:8 And, behold, thou hast with thee Shimei the son of

Gera, a Benjamite of Bahurim, which cursed me with a grievous curse in the day when I went to Mahanaim: but he came down to meet me at Jordan, and I sware to him by the LORD, saying, I will not put thee to death with the sword.

filium Gera filii Iemini de Baurim qui maledixit mihi maledictione pessima quando ibam ad Castra sed quia descendit mihi in occursum cum transirem Iordanem et iuraui ei per Dominum dicens non te interficiam gladio 2:9 tu noli pati 2:9 Now therefore hold him not guiltless: for thou art a

wise man, and knowest what thou oughtest to do unto him; but his hoar head bring thou down to the grave with blood.

esse eum innoxium uir autem sapiens es et scies quae facias

ei deducesque canos eius cum sanguine ad infernum 2:10 2:10 So David slept with his fathers, and was buried in the

city of David.

dormiuit igitur Dauid cum patribus suis et sepultus est in ciuitate Dauid 2:11 dies autem quibus regnauit Dauid su- 2:11 And the days that David reigned over Israel were

forty years: seven years reigned he in Hebron, and thirty and three years reigned he in Jerusalem.

per Israhel quadraginta anni sunt in Hebron regnauit septem annis in Hierusalem triginta tribus 2:12 Salomon autem se- 2:12 Then sat Solomon upon the throne of David his fa-

ther; and his kingdom was established greatly.

dit super thronum Dauid patris sui et firmatum est regnum eius nimis 2:13 et ingressus est Adonias filius Aggith ad 2:13 And Adonijah the son of Haggith came to Bathsheba

the mother of Solomon. And she said, Comest thou peaceably? And he said, Peaceably.

Bethsabee matrem Salomonis quae dixit ei pacificusne ingressus tuus qui respondit pacificus 2:14 addiditque sermo 2:14 He said moreover, I have somewhat to say unto thee.

And she said, Say on.

mihi est ad te cui ait loquere et ille 2:15 tu inquit nosti quia 2:15 And he said, Thou knowest that the kingdom was

mine, and that all Israel set their faces on me, that I should reign: howbeit the kingdom is turned about, and is become

meum erat regnum et me proposuerat omnis Israhel sibi in my brother’s: for it was his from the LORD. regem sed translatum est regnum et factum est fratris mei a Domino enim constitutum est ei 2:16 nunc ergo petitionem 2:16 And now I ask one petition of thee, deny me not. And

she said unto him, Say on.

unam deprecor a te ne confundas faciem meam quae dixit ad eum loquere 2:17 et ille ait precor ut dicas Salomoni regi 2:17 And he said, Speak, I pray thee, unto Solomon the

king, (for he will not say thee nay,) that he give me Abishag the Shunammite to wife.

neque enim negare tibi quicquam potest ut det mihi Abisag

Sunamitin uxorem 2:18 et ait Bethsabee bene ego loquar 2:18 And Bathsheba said, Well; I will speak for thee unto

the king.

pro te regi 2:19 uenit ergo Bethsabee ad regem Salomonem 2:19 Bathsheba therefore went unto king Solomon, to

speak unto him for Adonijah. And the king rose up to meet her, and bowed himself unto her, and sat down on his throne, and caused a seat to be set for the king’s mother; and she sat on his right hand.

ut loqueretur ei pro Adonia et surrexit rex in occursum eius adorauitque eam et sedit super thronum suum positus quoque est thronus matri regis quae sedit ad dexteram eius 2:20 2:20 Then she said, I desire one small petition of thee; I

pray thee, say me not nay. And the king said unto her, Ask on, my mother: for I will not say thee nay.

Biblia Sacra Vulgata

dixitque ei petitionem unam paruulam ego deprecor a te ne confundas faciem meam dixit ei rex pete mater mi neque

2:21 And she said, Let Abishag the Shunammite be given enim fas est ut auertam faciem tuam 2:21 quae ait detur to Adonijah thy brother to wife.

2:22 And king Solomon answered and said unto his Abisag Sunamitis Adoniae fratri tuo uxor 2:22 respondit-

mother, And why dost thou ask Abishag the Shunammite for Adonijah? ask for him the kingdom also; for he is mine elder brother; even for him, and for Abiathar the priest, and for Joab the son of Zeruiah.

que rex Salomon et dixit matri suae quare postulas Abisag Sunamitin Adoniae postula ei et regnum ipse est enim frater meus maior me et habet Abiathar sacerdotem et Ioab filium

2:23 Then king Solomon sware by the LORD, saying, God Saruiae 2:23 iurauit itaque rex Salomon per Dominum di-

do so to me, and more also, if Adonijah have not spoken this word against his own life.

cens haec faciat mihi Deus et haec addat quia contra animam

2:24 Now therefore, as the LORD liveth, which hath es- suam locutus est Adonias uerbum hoc 2:24 et nunc uiuit tablished me, and set me on the throne of David my father, and who hath made me an house, as he promised, Adoni-

jah shall be put to death this day. Dominus qui firmauit me et conlocauit super solium Dauid

patris mei et qui fecit mihi domum sicut locutus est quia ho-

2:25 And king Solomon sent by the hand of Benaiah the die occidetur Adonias 2:25 misitque rex Salomon per ma-

son of Jehoiada; and he fell upon him that he died.

num Banaiae filii Ioiadae qui interfecit eum et mortuus est

2:26 And unto Abiathar the priest said the king, Get thee 2:26 Abiathar quoque sacerdoti dixit rex uade in Anathot ad

to Anathoth, unto thine own fields; for thou art worthy of death: but I will not at this time put thee to death, because thou barest the ark of the LORD God before David my father, and because thou hast been afflicted in all wherein my father was afflicted.

agrum tuum es quidem uir mortis sed hodie te non interficiam quia portasti arcam Domini Dei coram Dauid patre meo et sustinuisti laborem in omnibus in quibus laborauit

2:27 So Solomon thrust out Abiathar from being priest pater meus 2:27 eiecit ergo Salomon Abiathar ut non esset

unto the LORD; that he might fulfil the word of the LORD, which he spake concerning the house of Eli in Shiloh.

sacerdos Domini ut impleretur sermo Domini quem locutus

2:28 Then tidings came to Joab: for Joab had turned after est super domum Heli in Silo 2:28 uenit autem nuntius ad Adonijah, though he turned not after Absalom. And Joab fled unto the tabernacle of the LORD, and caught hold on

the horns of the altar. Ioab quod Ioab declinasset post Adoniam et post Absalom

non declinasset fugit ergo Ioab in tabernaculum Domini et

2:29 And it was told king Solomon that Joab was fled unto adprehendit cornu altaris 2:29 nuntiatumque est regi Salo-

the tabernacle of the LORD; and, behold, he is by the altar. Then Solomon sent Benaiah the son of Jehoiada, saying,

Go, fall upon him. moni quod fugisset Ioab in tabernaculum Domini et esset

iuxta altare misitque Salomon Banaiam filium Ioiadae di-

2:30 And Benaiah came to the tabernacle of the LORD, cens uade interfice eum 2:30 uenit Banaias ad tabernacu-

and said unto him, Thus saith the king, Come forth. And he said, Nay; but I will die here. And Benaiah brought the king word again, saying, Thus said Joab, and thus he answered me.

lum Domini et dixit ei haec dicit rex egredere qui ait non egrediar sed hic moriar renuntiauit Banaias regi sermonem

2:31 And the king said unto him, Do as he hath said, and dicens haec locutus est Ioab et haec respondit mihi 2:31

fall upon him, and bury him; that thou mayest take away the innocent blood, which Joab shed, from me, and from

the house of my father. dixitque ei rex fac sicut locutus est et interfice eum et sepeli

et amouebis sanguinem innocentem qui effusus est a Ioab a

2:32 And the LORD shall return his blood upon his own me et a domo patris mei 2:32 et reddat Dominus sanguinem

head, who fell upon two men more righteous and better than he, and slew them with the sword, my father David not knowing thereof, to wit, Abner the son of Ner, captain of the host of Israel, and Amasa the son of Jether, captain of the host of Judah.

eius super caput eius quia interfecit duos uiros iustos melioresque se et occidit eos gladio patre meo Dauid ignorante Abner filium Ner principem militiae Israhel et Amasa filium Iether principem exercitus Iuda 2:33 et reuertetur sanguis 2:33 Their blood shall therefore return upon the head of

Joab, and upon the head of his seed for ever: but upon David, and upon his seed, and upon his house, and upon

illorum in caput Ioab et in caput seminis eius in sempiter- his throne, shall there be peace for ever from the LORD. num Dauid autem et semini eius et domui et throno illius sit pax usque in aeternum a Domino 2:34 ascendit itaque Ba- 2:34 So Benaiah the son of Jehoiada went up, and fell

upon him, and slew him: and he was buried in his own house in the wilderness.

naias filius Ioiadae et adgressus eum interfecit sepultusque est in domo sua in deserto 2:35 et constituit rex Banaiam 2:35 And the king put Benaiah the son of Jehoiada in his

room over the host: and Zadok the priest did the king put in the room of Abiathar.

filium Ioiadae pro eo super exercitum et Sadoc sacerdotem posuit pro Abiathar 2:36 misit quoque rex et uocauit Semei 2:36 And the king sent and called for Shimei, and said

unto him, Build thee an house in Jerusalem, and dwell there, and go not forth thence any whither.

dixitque ei aedifica tibi domum in Hierusalem et habita ibi et non egredieris inde huc atque illuc 2:37 quacumque au- 2:37 For it shall be, that on the day thou goest out, and

passest over the brook Kidron, thou shalt know for certain that thou shalt surely die: thy blood shall be upon thine

tem die egressus fueris et transieris torrentem Cedron scito own head. te interficiendum sanguis tuus erit super caput tuum 2:38 2:38 And Shimei said unto the king, The saying is good:

as my lord the king hath said, so will thy servant do. And Shimei dwelt in Jerusalem many days.

dixitque Semei regi bonus sermo sicut locutus est dominus meus rex sic faciet seruus tuus habitauit itaque Semei in Hierusalem diebus multis 2:39 factum est autem post an- 2:39 And it came to pass at the end of three years, that

two of the servants of Shimei ran away unto Achish son of Maachah king of Gath. And they told Shimei, saying,

nos tres ut fugerent serui Semei ad Achis filium Maacha re- Behold, thy servants be in Gath. gem Geth nuntiatumque est Semei quod serui eius essent in Geth 2:40 et surrexit Semei et strauit asinum suum iuitque 2:40 And Shimei arose, and saddled his ass, and went to

Gath to Achish to seek his servants: and Shimei went, and brought his servants from Gath.

in Geth ad Achis ad requirendos seruos suos et adduxit eos de Geth 2:41 nuntiatum est autem Salomoni quod isset Se- 2:41 And it was told Solomon that Shimei had gone from

Jerusalem to Gath, and was come again.

mei in Geth de Hierusalem et rediiset 2:42 et mittens uoca- 2:42 And the king sent and called for Shimei, and said

unto him, Did I not make thee to swear by the LORD, and protested unto thee, saying, Know for a certain, on the day thou goest out, and walkest abroad any whither, that thou shalt surely die? and thou saidst unto me, The word that I have heard is good.

uit eum dixitque illi nonne testificatus sum tibi per Dominum et praedixi tibi quacumque die egressus ieris huc et illuc scito te esse moriturum et respondisti mihi bonus sermo audiui 2:43 quare ergo non custodisti iusiurandum Domini et prae- 2:43 Why then hast thou not kept the oath of the LORD,

and the commandment that I have charged thee with?

ceptum quod praeceperam tibi 2:44 dixitque rex ad Semei 2:44 The king said moreover to Shimei, Thou knowest

all the wickedness which thine heart is privy to, that thou didst to David my father: therefore the LORD shall return

tu nosti omne malum cuius tibi conscium est cor tuum quod thy wickedness upon thine own head; fecisti Dauid patri meo reddidit Dominus malitiam tuam in caput tuum 2:45 et rex Salomon benedictus et thronus Da- 2:45 And king Solomon shall be blessed, and the throne

of David shall be established before the LORD for ever.

uid erit stabilis coram Domino usque in sempiternum 2:46 2:46 So the king commanded Benaiah the son of Jehoiada;

which went out, and fell upon him, that he died. And the kingdom was established in the hand of Solomon.

Biblia Sacra Vulgata

iussit itaque rex Banaiae filio Ioiadae qui egressus percussit eum et mortuus est

3:1 And Solomon made affinity with Pharaoh king of 3:1 confirmatum est igitur regnum in manu Salomonis et

Egypt, and took Pharaoh’s daughter, and brought her into the city of David, until he had made an end of building his own house, and the house of the LORD, and the wall of Jerusalem round about.

adfinitate coniunctus est Pharaoni regi Aegypti accepit namque filiam eius et adduxit in ciuitatem Dauid donec conpleret aedificans domum suam et domum Domini et murum Hie-

3:2 Only the people sacrificed in high places, because rusalem per circuitum 3:2 et tamen populus immolabat in

there was no house built unto the name of the LORD, until those days.

excelsis non enim aedificatum erat templum nomini Domini

3:3 And Solomon loved the LORD, walking in the statutes usque in die illo 3:3 dilexit autem Salomon Dominum amof David his father: only he sacrificed and burnt incense in high places.

bulans in praeceptis Dauid patris sui excepto quod in excel-

3:4 And the king went to Gibeon to sacrifice there; for sis immolabat et accendebat thymiama 3:4 abiit itaque in

that was the great high place: a thousand burnt offerings did Solomon offer upon that altar.

Gabaon ut immolaret ibi illud quippe erat excelsum maximum mille hostias in holocaustum obtulit Salomon super al-

3:5 In Gibeon the LORD appeared to Solomon in a dream tare illud in Gabaon 3:5 apparuit Dominus Salomoni per

by night: and God said, Ask what I shall give thee.

3:6 And Solomon said, Thou hast shewed unto thy ser- somnium nocte dicens postula quod uis ut dem tibi 3:6 et

vant David my father great mercy, according as he walked before thee in truth, and in righteousness, and in uprightness of heart with thee; and thou hast kept for him this great kindness, that thou hast given him a son to sit on his throne, as it is this day.

ait Salomon tu fecisti cum seruo tuo Dauid patre meo misericordiam magnam sicut ambulauit in conspectu tuo in ueritate et iustitia et recto corde tecum custodisti ei misericordiam tuam grandem et dedisti ei filium sedentem super thronum

3:7 And now, O LORD my God, thou hast made thy ser- eius sicut et hodie 3:7 et nunc Domine Deus tu regnare fe-

vant king instead of David my father: and I am but a little child: I know not how to go out or come in.

cisti seruum tuum pro Dauid patre meo ego autem sum puer

3:8 And thy servant is in the midst of thy people which paruus et ignorans egressum et introitum meum 3:8 et ser-

thou hast chosen, a great people, that cannot be numbered nor counted for multitude.

uus tuus in medio est populi quem elegisti populi infiniti qui

3:9 Give therefore thy servant an understanding heart to numerari et supputari non potest prae multitudine 3:9 dabis

judge thy people, that I may discern between good and bad: for who is able to judge this thy so great a people?

ergo seruo tuo cor docile ut iudicare possit populum tuum et discernere inter malum et bonum quis enim potest iudicare

3:10 And the speech pleased the LORD, that Solomon had populum istum populum tuum hunc multum 3:10 placuit asked this thing.

ergo sermo coram Domino quod Salomon rem huiuscemodi

3:11 And God said unto him, Because thou hast asked this postulasset 3:11 et dixit Deus Salomoni quia postulasti uer-

thing, and hast not asked for thyself long life; neither hast asked riches for thyself, nor hast asked the life of thine enemies; but hast asked for thyself understanding to discern judgment;

bum hoc et non petisti tibi dies multos nec diuitias aut animam inimicorum tuorum sed postulasti tibi sapientiam ad

3:12 Behold, I have done according to thy words: lo, I discernendum iudicium 3:12 ecce feci tibi secundum ser-

have given thee a wise and an understanding heart; so that there was none like thee before thee, neither after thee

shall any arise like unto thee. mones tuos et dedi tibi cor sapiens et intellegens in tantum ut nullus ante te similis tui fuerit nec post te surrecturus sit 3:13 sed et haec quae non postulasti dedi tibi diuitias scili- 3:13 And I have also given thee that which thou hast not

asked, both riches, and honour: so that there shall not be any among the kings like unto thee all thy days.

cet et gloriam ut nemo fuerit similis tui in regibus cunctis retro diebus 3:14 si autem ambulaueris in uiis meis et cus- 3:14 And if thou wilt walk in my ways, to keep my statutes

and my commandments, as thy father David did walk, then I will lengthen thy days.

todieris praecepta mea et mandata mea sicut ambulauit pater tuus longos faciam dies tuos 3:15 igitur euigilauit Salomon 3:15 And Solomon awoke; and, behold, it was a dream.

And he came to Jerusalem, and stood before the ark of the covenant of the LORD, and offered up burnt offerings, and offered peace offerings, and made a feast to all his servants.

et intellexit quod esset somnium cumque uenisset Hierusalem stetit coram arca foederis Domini et obtulit holocausta et fecit uictimas pacificas et grande conuiuium uniuersis famulis suis 3:16 tunc uenerunt duae mulieres meretrices ad 3:16 Then came there two women, that were harlots, unto

the king, and stood before him.

regem steteruntque coram eo 3:17 quarum una ait obsecro 3:17 And the one woman said, O my lord, I and this

woman dwell in one house; and I was delivered of a child with her in the house.

mi domine ego et mulier haec habitabamus in domo una et peperi apud eam in cubiculo 3:18 tertia uero die postquam 3:18 And it came to pass the third day after that I was de-

livered, that this woman was delivered also: and we were together; there was no stranger with us in the house, save

ego peperi peperit et haec et eramus simul nullusque alius we two in the house. in domo nobiscum exceptis nobis duabus 3:19 mortuus est 3:19 And this woman’s child died in the night; because

she overlaid it.

autem filius mulieris huius nocte dormiens quippe oppressit eum 3:20 et consurgens intempesta nocte silentio tulit 3:20 And she arose at midnight, and took my son from

beside me, while thine handmaid slept, and laid it in her bosom, and laid her dead child in my bosom.

filium meum de latere meo ancillae tuae dormientis et conlocauit in sinu suo suum autem filium qui erat mortuus posuit in sinu meo 3:21 cumque surrexissem mane ut darem 3:21 And when I rose in the morning to give my child

suck, behold, it was dead: but when I had considered it in the morning, behold, it was not my son, which I did bear.

lac filio meo apparuit mortuus quem diligentius intuens clara luce deprehendi non esse meum quem genueram 3:22 res- 3:22 And the other woman said, Nay; but the living is my

son, and the dead is thy son. And this said, No; but the dead is thy son, and the living is my son. Thus they spake

ponditque altera mulier non est ita sed filius tuus mortuus before the king. est meus autem uiuit e contrario illa dicebat mentiris filius quippe meus uiuit et filius tuus mortuus est atque in hunc modum contendebant coram rege 3:23 tunc rex ait haec di- 3:23 Then said the king, The one saith, This is my son that

liveth, and thy son is the dead: and the other saith, Nay; but thy son is the dead, and my son is the living.

cit filius meus uiuit et filius tuus mortuus est et ista respondit non sed filius tuus mortuus est et filius meus uiuit 3:24 dixit 3:24 And the king said, Bring me a sword. And they

brought a sword before the king.

ergo rex adferte mihi gladium cumque adtulissent gladium coram rege 3:25 diuidite inquit infantem uiuum in duas par- 3:25 And the king said, Divide the living child in two, and

give half to the one, and half to the other.

tes et date dimidiam partem uni et dimidiam partem alteri 3:26 dixit autem mulier cuius filius erat uiuus ad regem com- 3:26 Then spake the woman whose the living child was

unto the king, for her bowels yearned upon her son, and she said, O my lord, give her the living child, and in no wise slay it. But the other said, Let it be neither mine nor thine, but divide it.

mota sunt quippe uiscera eius super filio suo obsecro domine

Biblia Sacra Vulgata

date illi infantem uiuum et nolite interficere eum contra illa

3:27 Then the king answered and said, Give her the living dicebat nec mihi nec tibi sit diuidatur 3:27 respondens rex

child, and in no wise slay it: she is the mother thereof.

ait date huic infantem uiuum et non occidatur haec est mater

3:28 And all Israel heard of the judgment which the king eius 3:28 audiuit itaque omnis Israhel iudicium quod iudi-

had judged; and they feared the king: for they saw that the wisdom of God was in him, to do judgment.

casset rex et timuerunt regem uidentes sapientiam Dei esse in eo ad faciendum iudicium

4:1 So king Solomon was king over all Israel. 4:1 erat autem rex Salomon regnans super omnem Isra-4:2 And these were the princes which he had; Azariah the hel 4:2 et hii principes quos habebat Azarias filius Sadoc

son of Zadok the priest,

4:3 Elihoreph and Ahiah, the sons of Shisha, scribes; Je- sacerdos 4:3 Helioreph et Ahia filii Sesa scribae Iosaphat

hoshaphat the son of Ahilud, the recorder.

4:4 And Benaiah the son of Jehoiada was over the host: filius Ahilud a commentariis 4:4 Banaias filius Ioiadae su-

and Zadok and Abiathar were the priests:

4:5 And Azariah the son of Nathan was over the officers: per exercitum Sadoc autem et Abiathar sacerdotes 4:5 Aza-

and Zabud the son of Nathan was principal officer, and the king’s friend:

rias filius Nathan super eos qui adsistebant regi Zabud fi-

4:6 And Ahishar was over the household: and Adoniram lius Nathan sacerdos amicus regis 4:6 et Ahisar praeposi-

the son of Abda was over the tribute. 4:7 And Solomon had twelve officers over all Israel, which tus domus et Adoniram filius Abda super tributa 4:7 habeprovided victuals for the king and his household: each man his month in a year made provision.

bat autem Salomon duodecim praefectos super omnem Israhel qui praebebant annonam regi et domui eius per singulos enim menses in anno singuli necessaria ministrabant

4:8 And these are their names: The son of Hur, in mount 4:8 et haec nomina eorum Benhur in monte Ephraim 4:9 Ephraim: 4:9 The son of Dekar, in Makaz, and in Shaalbim, and Bethshemesh, and Elonbethhanan:

Bendecar in Macces et in Salebbim et in Bethsemes et He-

4:10 The son of Hesed, in Aruboth; to him pertained So- lon Bethanan 4:10 Benesed in Araboth ipsius erat Soccho

choh, and all the land of Hepher: 4:11 The son of Abinadab, in all the region of Dor; which et omnis terra Epher 4:11 Benabinadab cuius omnis Nephad Taphath the daughter of Solomon to wife: 4:12 Baana the son of Ahilud; to him pertained Taanach thad Dor Tapheth filiam Salomonis habebat uxorem 4:12 and Megiddo, and all Bethshean, which is by Zartanah beneath Jezreel, from Bethshean to Abelmeholah, even unto

the place that is beyond Jokneam: Bana filius Ahilud regebat Thanac et Mageddo et uniuersam

Bethsan quae est iuxta Sarthana subter Hiezrahel a Bethsan

4:13 The son of Geber, in Ramothgilead; to him pertained usque Abelmeula e regione Iecmaan 4:13 Bengaber in Rathe towns of Jair the son of Manasseh, which are in Gilead; to him also pertained the region of Argob, which is in Bashan, threescore great cities with walls and brasen bars: moth Galaad habebat Auothiair filii Manasse in Galaad ipse

praeerat in omni regione Argob quae est in Basan sexaginta ciuitatibus magnis atque muratis quae habebant seras aereas

4:14 Ahinadab the son of Iddo had Mahanaim: 4:14 Ahinadab filius Addo praeerat in Manaim 4:15 Ahi-

4:15 Ahimaaz was in Naphtali; he also took Basmath the daughter of Solomon to wife:

maas in Nepthali sed et ipse habebat Basmath filiam Salo-

4:16 Baanah the son of Hushai was in Asher and in Aloth: monis in coniugio 4:16 Baana filius Vsi in Aser et in Ba-

4:17 Jehoshaphat the son of Paruah, in Issachar: lod 4:17 Iosaphat filius Pharue in Isachar 4:18 Semei fi-

4:18 Shimei the son of Elah, in Benjamin:

4:19 Geber the son of Uri was in the country of Gilead, in lius Hela in Beniamin 4:19 Gaber filius Vri in terra Galaad

the country of Sihon king of the Amorites, and of Og king of Bashan; and he was the only officer which was in the land.

in terra Seon regis Amorrei et Og regis Basan super omnia quae erant in illa terra 4:20 Iuda et Israhel innumerabi- 4:20 Judah and Israel were many, as the sand which is

by the sea in multitude, eating and drinking, and making merry.

les sicut harena maris in multitudine comedentes et bibentes atque laetantes 4:21 Salomon autem erat in dicione sua 4:21 And Solomon reigned over all kingdoms from the

river unto the land of the Philistines, and unto the border of Egypt: they brought presents, and served Solomon all

habens omnia regna sicut a flumine terrae Philisthim usque the days of his life. ad terminum Aegypti offerentium sibi munera et seruientium ei cunctis diebus uitae eius 4:22 erat autem cibus Sa- 4:22 And Solomon’s provision for one day was thirty mea-

sures of fine flour, and threescore measures of meal,

lomonis per dies singulos triginta chori similae et sexaginta chori farinae 4:23 decem boues pingues et uiginti boues pa- 4:23 Ten fat oxen, and twenty oxen out of the pastures,

and an hundred sheep, beside harts, and roebucks, and fallowdeer, and fatted fowl.

scuales et centum arietes excepta uenatione ceruorum caprearum atque bubalorum et auium altilium 4:24 ipse enim 4:24 For he had dominion over all the region on this side

the river, from Tiphsah even to Azzah, over all the kings on this side the river: and he had peace on all sides round

obtinebat omnem regionem quae erat trans flumen quasi a about him. Thapsa usque Gazam et cunctos reges illarum regionum et habebat pacem ex omni parte in circuitu 4:25 habitabatque 4:25 And Judah and Israel dwelt safely, every man under

his vine and under his fig tree, from Dan even to Beersheba, all the days of Solomon.

Iudas et Israhel absque timore ullo unusquisque sub uite sua et sub ficu sua a Dan usque Bersabee cunctis diebus Salomonis 4:26 et habebat Salomon quadraginta milia praese- 4:26 And Solomon had forty thousand stalls of horses for

his chariots, and twelve thousand horsemen.

pia equorum currulium et duodecim milia equestrium 4:27 4:27 And those officers provided victual for king

Solomon, and for all that came unto king Solomon’s table, every man in his month: they lacked nothing.

nutriebantque eos supradicti regis praefecti sed et necessaria mensae regis Salomonis cum ingenti cura praebebant in

tempore suo 4:28 hordeum quoque et paleas equorum et iu- 4:28 Barley also and straw for the horses and dromedaries brought they unto the place where the officers were, every man according to his charge.

mentorum deferebant in locum ubi erat rex iuxta constitutum sibi 4:29 dedit quoque Deus sapientiam Salomoni et 4:29 And God gave Solomon wisdom and understanding

exceeding much, and largeness of heart, even as the sand that is on the sea shore.

prudentiam multam nimis et latitudinem cordis quasi harenam quae est in litore maris 4:30 et praecedebat sapientia 4:30 And Solomon’s wisdom excelled the wisdom of all

the children of the east country, and all the wisdom of Egypt.

Salomonis sapientiam omnium Orientalium et Aegyptiorum 4:31 et erat sapientior cunctis hominibus sapientior Aethan 4:31 For he was wiser than all men; than Ethan the

Ezrahite, and Heman, and Chalcol, and Darda, the sons of Mahol: and his fame was in all nations round about.

Ezraita et Heman et Chalcal et Dorda filiis Maol et erat nominatus in uniuersis gentibus per circuitum 4:32 locutus est 4:32 And he spake three thousand proverbs: and his songs

were a thousand and five.

quoque Salomon tria milia parabolas et fuerunt carmina eius quinque et mille 4:33 et disputauit super lignis a cedro quae 4:33 And he spake of trees, from the cedar tree that is in

Lebanon even unto the hyssop that springeth out of the wall: he spake also of beasts, and of fowl, and of creeping

est in Libano usque ad hysopum quae egreditur de pariete et things, and of fishes. disseruit de iumentis et uolucribus et reptilibus et piscibus

Biblia Sacra Vulgata

4:34 And there came of all people to hear the wisdom of 4:34 et ueniebant de cunctis populis ad audiendam sapien-

Solomon, from all kings of the earth, which had heard of his wisdom.

tiam Salomonis et ab uniuersis regibus terrae qui audiebant sapientiam eius

5:1 And Hiram king of Tyre sent his servants unto 5:1 misit quoque Hiram rex Tyri seruos suos ad Salomo-

Solomon; for he had heard that they had anointed him king in the room of his father: for Hiram was ever a lover of

David. nem audiuit enim quod ipsum unxissent regem pro patre eius 5:2 And Solomon sent to Hiram, saying, quia amicus fuerat Hiram Dauid omni tempore 5:2 misit 5:3 Thou knowest how that David my father could not autem et Salomon ad Hiram dicens 5:3 tu scis uoluntatem

build an house unto the name of the LORD his God for the wars which were about him on every side, until the

LORD put them under the soles of his feet. Dauid patris mei et quia non potuerit aedificare domum no-

mini Domini Dei sui propter bella inminentia per circuitum

5:4 But now the LORD my God hath given me rest on donec daret Dominus eos sub uestigio pedum eius 5:4 nunc

every side, so that there is neither adversary nor evil occurrent.

autem requiem dedit Deus meus mihi per circuitum non est

5:5 And, behold, I purpose to build an house unto the name Satan neque occursus malus 5:5 quam ob rem cogito aediof the LORD my God, as the LORD spake unto David my father, saying, Thy son, whom I will set upon thy throne

in thy room, he shall build an house unto my name. ficare templum nomini Domini Dei mei sicut locutus est Do-

minus Dauid patri meo dicens filius tuus quem dabo pro te

5:6 Now therefore command thou that they hew me cedar super solium tuum ipse aedificabit domum nomini meo 5:6

trees out of Lebanon; and my servants shall be with thy servants: and unto thee will I give hire for thy servants according to all that thou shalt appoint: for thou knowest that there is not among us any that can skill to hew timber like unto the Sidonians.

praecipe igitur ut praecidant mihi cedros de Libano et serui mei sint cum seruis tuis mercedem autem seruorum tuorum dabo tibi quamcumque praeceperis scis enim quoniam non est in populo meo uir qui nouerit ligna caedere sicut Sidonii

5:7 And it came to pass, when Hiram heard the words of 5:7 cum ergo audisset Hiram uerba Salomonis laetatus est

Solomon, that he rejoiced greatly, and said, Blessed be the LORD this day, which hath given unto David a wise son

over this great people. ualde et ait benedictus Dominus hodie qui dedit Dauid filium 5:8 And Hiram sent to Solomon, saying, I have consid- sapientissimum super populum hunc plurimum 5:8 et misit

ered the things which thou sentest to me for: and I will do all thy desire concerning timber of cedar, and concerning

timber of fir. Hiram ad Salomonem dicens audiui quaecumque mandasti

mihi ego faciam omnem uoluntatem tuam in lignis cedrinis

5:9 My servants shall bring them down from Lebanon unto et abiegnis 5:9 serui mei deponent ea de Libano ad mare et

the sea: and I will convey them by sea in floats unto the place that thou shalt appoint me, and will cause them to be discharged there, and thou shalt receive them: and thou shalt accomplish my desire, in giving food for my household.

ego conponam ea in ratibus in mari usque ad locum quem significaueris mihi et adplicabo ea ibi et tu tolles ea prae-

5:10 So Hiram gave Solomon cedar trees and fir trees ac- bebisque necessaria mihi ut detur cibus domui meae 5:10

cording to all his desire.

itaque Hiram dabat Salomoni ligna cedrina et ligna abiegna

5:11 And Solomon gave Hiram twenty thousand measures iuxta omnem uoluntatem eius 5:11 Salomon autem praebe-

of wheat for food to his household, and twenty measures of pure oil: thus gave Solomon to Hiram year by year.

bat Hiram uiginti milia chororum tritici in cibum domui eius et uiginti choros purissimi olei haec tribuebat Salomon Hi-

5:12 And the LORD gave Solomon wisdom, as he ram per annos singulos 5:12 dedit quoque Dominus sapien-

promised him: and there was peace between Hiram and Solomon; and they two made a league together.

tiam Salomoni sicut locutus est ei et erat pax inter Hiram et Salomonem et percusserunt foedus ambo 5:13 legitque rex 5:13 And king Solomon raised a levy out of all Israel; and

the levy was thirty thousand men.

Salomon operas de omni Israhel et erat indictio triginta milia uirorum 5:14 mittebatque eos in Libanum decem milia 5:14 And he sent them to Lebanon, ten thousand a month

by courses: a month they were in Lebanon, and two months at home: and Adoniram was over the levy.

per menses singulos uicissim ita ut duobus mensibus essent in domibus suis et Adoniram erat super huiuscemodi indictione 5:15 fuerunt itaque Salomoni septuaginta milia eorum 5:15 And Solomon had threescore and ten thousand that

bare burdens, and fourscore thousand hewers in the mountains;

qui onera portabant et octoginta milia latomorum in monte 5:16 absque praepositis qui praeerant singulis operibus nu- 5:16 Beside the chief of Solomon’s officers which were

over the work, three thousand and three hundred, which ruled over the people that wrought in the work.

mero trium milium et trecentorum praecipientium populo et his qui faciebant opus 5:17 praecepitque rex ut tollerent la- 5:17 And the king commanded, and they brought great

stones, costly stones, and hewed stones, to lay the foundation of the house.

pides grandes lapides pretiosos in fundamentum templi et quadrarent eos 5:18 quos dolauerunt cementarii Salomonis 5:18 And Solomon’s builders and Hiram’s builders did

hew them, and the stonesquarers: so they prepared timber and stones to build the house.

et cementarii Hiram porro Biblii praeparauerunt ligna et lapides ad aedificandam domum

6:1 factum est igitur quadringentesimo et octogesimo anno 6:1 And it came to pass in the four hundred and eightieth

year after the children of Israel were come out of the land of Egypt, in the fourth year of Solomon’s reign over Israel, in the month Zif, which is the second month, that he began to build the house of the LORD.

egressionis filiorum Israhel de terra Aegypti in anno quarto mense zio ipse est mensis secundus regis Salomonis super Israhel aedificare coepit domum Domino 6:2 domus autem 6:2 And the house which king Solomon built for the

LORD, the length thereof was threescore cubits, and the breadth thereof twenty cubits, and the height thereof thirty

quam aedificabat rex Salomon Domino habebat sexaginta cubits. cubitos in longitudine et uiginti cubitos in latitudine et triginta cubitos in altitudine 6:3 et porticus erat ante temp- 6:3 And the porch before the temple of the house, twenty

cubits was the length thereof, according to the breadth of the house; and ten cubits was the breadth thereof before

lum uiginti cubitorum longitudinis iuxta mensuram latitudi- the house. nis templi et habebat decem cubitos latitudinis ante faciem templi 6:4 fecitque in templo fenestras obliquas 6:5 et ae- 6:4 And for the house he made windows of narrow lights.

6:5 And against the wall of the house he built chambers round about, against the walls of the house round about, both of the temple and of the oracle: and he made chambers round about:

dificauit super parietem templi tabulata per gyrum in parietibus domus per circuitum templi et oraculi et fecit latera in circuitu 6:6 tabulatum quod subter erat quinque cubitos 6:6 The nethermost chamber was five cubits broad, and

the middle was six cubits broad, and the third was seven cubits broad: for without in the wall of the house he made narrowed rests round about, that the beams should not be fastened in the walls of the house.

habebat latitudinis et medium tabulatum sex cubitorum latitudinis et tertium tabulatum septem habens cubitos latitudinis trabes autem posuit in domo per circuitum forinsecus ut non hererent muris templi 6:7 domus autem cum aedi- 6:7 And the house, when it was in building, was built of

stone made ready before it was brought thither: so that there was neither hammer nor axe nor any tool of iron

ficaretur lapidibus dedolatis atque perfectis aedificata est et heard in the house, while it was in building.

Biblia Sacra Vulgata

malleus et securis et omne ferramentum non sunt audita in

6:8 The door for the middle chamber was in the right side domo cum aedificaretur 6:8 ostium lateris medii in parte

of the house: and they went up with winding stairs into the middle chamber, and out of the middle into the third.

erat domus dexterae et per cocleam ascendebant in medium

6:9 So he built the house, and finished it; and covered the cenaculum et a medio in tertium 6:9 et aedificauit domum house with beams and boards of cedar.

et consummauit eam texit quoque domum laquearibus ce-

6:10 And then he built chambers against all the house, five drinis 6:10 et aedificauit tabulatum super omnem domum cubits high: and they rested on the house with timber of cedar.

quinque cubitis altitudinis et operuit domum lignis cedri-

6:11 And the word of the LORD came to Solomon, saying, nis 6:11 et factus est sermo Domini ad Salomonem dicens

6:12 Concerning this house which thou art in building, if 6:12 domus haec quam aedificas si ambulaueris in praecepthou wilt walk in my statutes, and execute my judgments, and keep all my commandments to walk in them; then will I perform my word with thee, which I spake unto David thy father:

tis meis et iudicia mea feceris et custodieris omnia mandata mea gradiens per ea firmabo sermonem meum tibi quem lo-

6:13 And I will dwell among the children of Israel, and cutus sum ad Dauid patrem tuum 6:13 et habitabo in mewill not forsake my people Israel.

dio filiorum Israhel et non derelinquam populum meum Isra-

6:14 So Solomon built the house, and finished it. hel 6:14 igitur aedificauit Salomon domum et consummauit 6:15 And he built the walls of the house within with boards eam 6:15 et aedificauit parietes domus intrinsecus tabula-

of cedar, both the floor of the house, and the walls of the ceiling: and he covered them on the inside with wood, and

covered the floor of the house with planks of fir. tis cedrinis a pauimento domus usque ad summitatem parie-

tum et usque ad laquearia operuit lignis intrinsecus et texit

6:16 And he built twenty cubits on the sides of the house, pauimentum domus tabulis abiegnis 6:16 aedificauitque ui-

both the floor and the walls with boards of cedar: he even built them for it within, even for the oracle, even for the

most holy place. ginti cubitorum ad posteriorem partem templi tabulata ced-

rina a pauimento usque ad superiora et fecit interiorem do-

6:17 And the house, that is, the temple before it, was forty mum oraculi in sanctum sanctorum 6:17 porro quadraginta cubits long.

6:18 And the cedar of the house within was carved with cubitorum erat ipsum templum pro foribus oraculi 6:18 et

knops and open flowers: all was cedar; there was no stone seen.

cedro omnis domus intrinsecus uestiebatur habens tornaturas suas et iuncturas fabrefactas et celaturas eminentes omnia cedrinis tabulis uestiebantur nec omnino lapis apparere

6:19 And the oracle he prepared in the house within, to set poterat in pariete 6:19 oraculum autem in medio domus in there the ark of the covenant of the LORD.

interiori parte fecerat ut poneret ibi arcam foederis Domini

6:20 And the oracle in the forepart was twenty cubits in 6:20 porro oraculum habebat uiginti cubitos longitudinis et

length, and twenty cubits in breadth, and twenty cubits in the height thereof: and he overlaid it with pure gold; and

so covered the altar which was of cedar. uiginti cubitos latitudinis et uiginti cubitos altitudinis et ope-

ruit illud atque uestiuit auro purissimo sed et altare uesti-

6:21 So Solomon overlaid the house within with pure gold: uit cedro 6:21 domum quoque ante oraculum operuit auro and he made a partition by the chains of gold before the oracle; and he overlaid it with gold.

purissimo et adfixit lamminas clauis aureis 6:22 nihilque

6:22 And the whole house he overlaid with gold, until he had finished all the house: also the whole altar that was by the oracle he overlaid with gold.

erat in templo quod non auro tegeretur sed et totum altare oraculi texit auro 6:23 et fecit in oraculo duo cherubin de 6:23 And within the oracle he made two cherubims of

olive tree, each ten cubits high.

lignis oliuarum decem cubitorum altitudinis 6:24 quinque 6:24 And five cubits was the one wing of the cherub, and

five cubits the other wing of the cherub: from the uttermost part of the one wing unto the uttermost part of the other

cubitorum ala cherub una et quinque cubitorum ala cherub were ten cubits. altera id est decem cubitos habentes a summitate alae usque ad alae alterius summitatem 6:25 decem quoque cubitorum 6:25 And the other cherub was ten cubits: both the cheru-

bims were of one measure and one size.

erat cherub secundus mensura pari et opus unum erat in duobus cherubin 6:26 id est altitudinem habebat unus cherub 6:26 The height of the one cherub was ten cubits, and so

was it of the other cherub.

decem cubitorum et similiter cherub secundus 6:27 posuit- 6:27 And he set the cherubims within the inner house: and

they stretched forth the wings of the cherubims, so that the wing of the one touched the one wall, and the wing of the other cherub touched the other wall; and their wings touched one another in the midst of the house.

que cherubin in medio templi interioris extendebant autem alas suas cherubin et tangebat ala una parietem et ala cherub secundi tangebat parietem alterum alae autem alterae in media parte templi se inuicem contingebant 6:28 texit quo- 6:28 And he overlaid the cherubims with gold. que cherubin auro 6:29 et omnes parietes templi per circui- 6:29 And he carved all the walls of the house round about

with carved figures of cherubims and palm trees and open flowers, within and without.

tum scalpsit uariis celaturis et torno et fecit in eis cherubin et palmas et picturas uarias quasi prominentes de pariete et egredientes 6:30 sed et pauimentum domus texit auro in- 6:30 And the floors of the house he overlaid with gold,

within and without.

trinsecus et extrinsecus 6:31 et in ingressu oraculi fecit os- 6:31 And for the entering of the oracle he made doors of

olive tree: the lintel and side posts were a fifth part of the wall.

tiola de lignis oliuarum postesque angulorum quinque 6:32

6:32 The two doors also were of olive tree; and he carved upon them carvings of cherubims and palm trees and open flowers, and overlaid them with gold, and spread gold upon the cherubims, and upon the palm trees.

et duo ostia de lignis oliuarum et scalpsit in eis picturam cherubin et palmarum species et anaglyfa ualde prominentia et texit ea auro et operuit tam cherubin quam palmas et cetera auro 6:33 fecitque in introitum templi postes de lignis oli- 6:33 So also made he for the door of the temple posts of

olive tree, a fourth part of the wall.

uarum quadrangulatos 6:34 et duo ostia de lignis abiegnis 6:34 And the two doors were of fir tree: the two leaves of

the one door were folding, and the two leaves of the other door were folding.

altrinsecus et utrumque ostium duplex erat et se inuicem tenens aperiebatur 6:35 et scalpsit cherubin et palmas et ce- 6:35 And he carved thereon cherubims and palm trees and

open flowers: and covered them with gold fitted upon the carved work.

laturas ualde eminentes operuitque omnia lamminis aureis opere quadro ad regulam 6:36 et aedificauit atrium interius 6:36 And he built the inner court with three rows of hewed

stone, and a row of cedar beams.

tribus ordinibus lapidum politorum et uno ordine lignorum cedri 6:37 anno quarto fundata est domus Domini in mense 6:37 In the fourth year was the foundation of the house of

the LORD laid, in the month Zif:

zio 6:38 et in anno undecimo mense bul ipse est mensis oc- 6:38 And in the eleventh year, in the month Bul, which

is the eighth month, was the house finished throughout all the parts thereof, and according to all the fashion of it. So

tauus perfecta est domus in omni opere suo et in uniuersis was he seven years in building it. utensilibus aedificauitque eam annis septem

7:1 domum autem suam aedificauit Salomon tredecim an- 7:1 But Solomon was building his own house thirteen

years, and he finished all his house.

Biblia Sacra Vulgata

7:2 He built also the house of the forest of Lebanon; the nis et ad perfectum usque perduxit 7:2 aedificauit quoque

length thereof was an hundred cubits, and the breadth thereof fifty cubits, and the height thereof thirty cubits, upon four rows of cedar pillars, with cedar beams upon the pillars.

domum saltus Libani centum cubitorum longitudinis et quinquaginta cubitorum latitudinis et triginta cubitorum altitudinis et quattuor deambulacra inter columnas cedrinas ligna

7:3 And it was covered with cedar above upon the beams, quippe cedrina exciderat in columnas 7:3 et tabulatis ced-

that lay on forty five pillars, fifteen in a row.

rinis uestiuit totam cameram quae quadraginta quinque columnis sustentabatur unus autem ordo habebat columnas

7:4 And there were windows in three rows, and light was quindecim 7:4 contra se inuicem positas 7:5 et e regione

against light in three ranks. 7:5 And all the doors and posts were square, with the windows: and light was against light in three ranks.

se respicientes aequali spatio inter columnas et super co-

7:6 And he made a porch of pillars; the length thereof was lumnas quadrangulata ligna in cunctis aequalia 7:6 et por-

fifty cubits, and the breadth thereof thirty cubits: and the porch was before them: and the other pillars and the thick

beam were before them. ticum columnarum fecit quinquaginta cubitorum longitudinis et triginta cubitorum latitudinis et alteram porticum in facie maioris porticus et columnas et epistylia super colum-

7:7 Then he made a porch for the throne where he might nas 7:7 porticum quoque solii in qua tribunal est fecit et

judge, even the porch of judgment: and it was covered with cedar from one side of the floor to the other.

texit lignis cedrinis a pauimento usque ad summitatem 7:8

7:8 And his house where he dwelt had another court within the porch, which was of the like work. Solomon made also an house for Pharaoh’s daughter, whom he had taken to wife, like unto this porch.

et domuncula in qua sedetur ad iudicandum erat in media porticu simili opere domum quoque fecit filiae Pharaonis quam uxorem duxerat Salomon tali opere quali et hanc por-

7:9 All these were of costly stones, according to the mea- ticum 7:9 omnia lapidibus pretiosis qui ad normam quansures of hewed stones, sawed with saws, within and without, even from the foundation unto the coping, and so on

the outside toward the great court. dam atque mensuram tam intrinsecus quam extrinsecus ser-

rati erant a fundamento usque ad summitatem parietum et

7:10 And the foundation was of costly stones, even great intrinsecus usque ad atrium maius 7:10 fundamenta autem

stones, stones of ten cubits, and stones of eight cubits.

de lapidibus pretiosis lapidibus magnis decem siue octo cu-

7:11 And above were costly stones, after the measures of bitorum 7:11 et desuper lapides pretiosi aequalis mensurae

hewed stones, and cedars.

7:12 And the great court round about was with three rows secti erant similiterque de cedro 7:12 et atrium maius ro-

of hewed stones, and a row of cedar beams, both for the inner court of the house of the LORD, and for the porch

of the house. tundum trium ordinum de lapidibus sectis et unius ordinis

dolata cedro necnon et in atrio domus Domini interiori et

7:13 And king Solomon sent and fetched Hiram out of in porticu domus 7:13 misit quoque rex Salomon et tulit

Tyre.

7:14 He was a widow’s son of the tribe of Naphtali, and Hiram de Tyro 7:14 filium mulieris uiduae de tribu Nep-

his father was a man of Tyre, a worker in brass: and he was filled with wisdom, and understanding, and cunning to work all works in brass. And he came to king Solomon, and wrought all his work.

thali patre Tyrio artificem aerarium et plenum sapientia et intellegentia et doctrina ad faciendum omne opus ex aere qui cum uenisset ad regem Salomonem fecit omne opus eius

7:15 For he cast two pillars of brass, of eighteen cubits 7:15 et finxit duas columnas aereas decem et octo cubitohigh apiece: and a line of twelve cubits did compass either of them about.

rum altitudinis columnam unam et linea duodecim cubitorum ambiebat columnam utramque 7:16 duo quoque capi- 7:16 And he made two chapiters of molten brass, to set

upon the tops of the pillars: the height of the one chapiter was five cubits, and the height of the other chapiter was

tella fecit quae ponerentur super capita columnarum fusili five cubits: aere quinque cubitorum altitudinis capitellum unum et quinque cubitorum altitudinis capitellum alterum 7:17 et quasi 7:17 And nets of checker work, and wreaths of chain work,

for the chapiters which were upon the top of the pillars; seven for the one chapiter, and seven for the other chapiter.

in modum retis et catenarum sibi inuicem miro opere contextarum utrumque capitellum columnarum fusile erat septena uersuum retiacula in capitello uno et septena retiacula in capitello altero 7:18 et perfecit columnas et duos ordines per 7:18 And he made the pillars, and two rows round about

upon the one network, to cover the chapiters that were upon the top, with pomegranates: and so did he for the

circuitum retiaculorum singulorum ut tegerent capitella quae other chapiter. erant super summitatem malogranatorum eodem modo fecit et capitello secundo 7:19 capitella autem quae erant super 7:19 And the chapiters that were upon the top of the pillars

were of lily work in the porch, four cubits.

capita columnarum quasi opere lilii fabricata erant in porticu quattuor cubitorum 7:20 et rursum alia capitella in summi- 7:20 And the chapiters upon the two pillars had

pomegranates also above, over against the belly which was by the network: and the pomegranates were two hundred

tate columnarum desuper iuxta mensuram columnae contra in rows round about upon the other chapiter. retiacula malogranatorum autem ducenti ordines erant in circuitu capitelli secundi 7:21 et statuit duas columnas in por- 7:21 And he set up the pillars in the porch of the temple:

and he set up the right pillar, and called the name thereof Jachin: and he set up the left pillar, and called the name

ticum templi cumque statuisset columnam dexteram uocauit thereof Boaz. eam nomine Iachin similiter erexit columnam secundam et uocauit nomen eius Booz 7:22 et super capita columnarum 7:22 And upon the top of the pillars was lily work: so was

the work of the pillars finished.

opus in modum lilii posuit perfectumque est opus columnarum 7:23 fecit quoque mare fusile decem cubitorum a labio 7:23 And he made a molten sea, ten cubits from the one

brim to the other: it was round all about, and his height was five cubits: and a line of thirty cubits did compass it

usque ad labium rotundum in circuitu quinque cubitorum al- round about. titudo eius et resticula triginta cubitorum cingebat illud per circuitum 7:24 et scalptura subter labium circumibat illud 7:24 And under the brim of it round about there were

knops compassing it, ten in a cubit, compassing the sea round about: the knops were cast in two rows, when it

decem cubitis ambiens mare duo ordines scalpturarum hi- was cast. striatarum erant fusiles 7:25 et stabat super duodecim bo- 7:25 It stood upon twelve oxen, three looking toward the

north, and three looking toward the west, and three looking toward the south, and three looking toward the east: and the sea was set above upon them, and all their hinder parts were inward.

ues e quibus tres respiciebant ad aquilonem et tres ad occidentem et tres ad meridiem et tres ad orientem et mare super eos desuper erat quorum posteriora uniuersa intrinsecus latitabant 7:26 grossitudo autem luteris trium unciarum erat 7:26 And it was an hand breadth thick, and the brim

thereof was wrought like the brim of a cup, with flowers of lilies: it contained two thousand baths.

labiumque eius quasi labium calicis et folium repandi lilii duo milia batos capiebat 7:27 et fecit bases decem aereas 7:27 And he made ten bases of brass; four cubits was the

length of one base, and four cubits the breadth thereof, and three cubits the height of it.

Biblia Sacra Vulgata

quattuor cubitorum longitudinis bases singulas et quattuor

7:28 And the work of the bases was on this manner: they cubitorum latitudinis et trium cubitorum altitudinis 7:28 et

had borders, and the borders were between the ledges:

ipsum opus basium interrasile erat et scalpturae inter iunc-

7:29 And on the borders that were between the ledges were turas 7:29 et inter coronulas et plectas leones et boues et

lions, oxen, and cherubims: and upon the ledges there was a base above: and beneath the lions and oxen were certain

additions made of thin work. cherubin et in iuncturis similiter desuper et subter leones et 7:30 And every base had four brasen wheels, and plates boues quasi lora ex aere dependentia 7:30 et quattuor rotae

of brass: and the four corners thereof had undersetters: under the laver were undersetters molten, at the side of

every addition. per bases singulas et axes aerei et per quattuor partes quasi

umeruli subter luterem fusiles contra se inuicem respectan-

7:31 And the mouth of it within the chapiter and above was tes 7:31 os quoque luteris intrinsecus erat in capitis sum-

a cubit: but the mouth thereof was round after the work of the base, a cubit and an half: and also upon the mouth of it were gravings with their borders, foursquare, not round. mitate et quod forinsecus apparebat unius cubiti erat totum

rotundum pariterque habebat unum cubitum et dimidium in angulis autem columnarum uariae celaturae erant et media

7:32 And under the borders were four wheels; and the intercolumnia quadrata non rotunda 7:32 quattuor quoque

axletrees of the wheels were joined to the base: and the height of a wheel was a cubit and half a cubit.

rotae quae per quattuor angulos basis erant coherebant subter

7:33 And the work of the wheels was like the work of a basi una rota habebat altitudinis cubitum et semis 7:33 tales

chariot wheel: their axletrees, and their naves, and their felloes, and their spokes, were all molten.

autem rotae erant quales solent in curru fieri et axes earum et

7:34 And there were four undersetters to the four corners radii et canti et modioli omnia fusilia 7:34 nam et umeruli

of one base: and the undersetters were of the very base itself.

illi quattuor per singulos angulos basis unius ex ipsa basi fu-

7:35 And in the top of the base was there a round compass siles et coniuncti erant 7:35 in summitate autem basis erat

of half a cubit high: and on the top of the base the ledges thereof and the borders thereof were of the same.

quaedam rotunditas dimidii cubiti ita fabrefacta ut luter desuper possit inponi habens celaturas suas et scalpturas uarias

7:36 For on the plates of the ledges thereof, and on the bor- ex semet ipso 7:36 scalpsit quoque in tabulatis illis quae er-

ders thereof, he graved cherubims, lions, and palm trees, according to the proportion of every one, and additions

round about. ant ex aere et in angulis cherubin et leones et palmas quasi

in similitudinem stantis hominis ut non celata sed adposita

7:37 After this manner he made the ten bases: all of them per circuitum uiderentur 7:37 in hunc modum fecit decem

had one casting, one measure, and one size.

7:38 Then made he ten lavers of brass: one laver contained bases fusura una et mensura scalpturaque consimili 7:38 fe-

forty baths: and every laver was four cubits: and upon every one of the ten bases one laver.

cit quoque decem luteres aereos quadraginta batos capiebat luter unus eratque quattuor cubitorum singulosque luteres

7:39 And he put five bases on the right side of the house, per singulas id est decem bases posuit 7:39 et constituit de-

and five on the left side of the house: and he set the sea on the right side of the house eastward over against the south.

cem bases quinque ad dexteram partem templi et quinque ad sinistram mare autem posuit ad dexteram partem templi

7:40 And Hiram made the lavers, and the shovels, and the contra orientem ad meridiem 7:40 fecit ergo Hiram lebetas

basons. So Hiram made an end of doing all the work that he made king Solomon for the house of the LORD:

et scutras et amulas et perfecit omne opus regis Salomonis in templo Domini 7:41 columnas duas et funiculos capitu- 7:41 The two pillars, and the two bowls of the chapiters

that were on the top of the two pillars; and the two networks, to cover the two bowls of the chapiters which were

lorum super capitella columnarum duos et retiacula duo ut upon the top of the pillars; operirent duos funiculos qui erant super capita columnarum 7:42 et malogranata quadringenta in duobus retiaculis duos 7:42 And four hundred pomegranates for the two net-

works, even two rows of pomegranates for one network, to cover the two bowls of the chapiters that were upon the

uersus malogranatorum in retiaculis singulis ad operiendos pillars; funiculos capitellorum qui erant super capita columnarum 7:43 et bases decem et luteres decem super bases 7:44 et 7:43 And the ten bases, and ten lavers on the bases;

7:44 And one sea, and twelve oxen under the sea;

mare unum et boues duodecim subter mare 7:45 et lebetas 7:45 And the pots, and the shovels, and the basons: and all these vessels, which Hiram made to king Solomon for the house of the LORD, were of bright brass.

et scutras et amulas omnia uasa quae fecit Hiram regi Salomoni in domo Domini de aurichalco erant 7:46 in campestri 7:46 In the plain of Jordan did the king cast them, in the

clay ground between Succoth and Zarthan.

regione Iordanis fudit ea rex in argillosa terra inter Socchoth et Sarthan 7:47 et posuit Salomon omnia uasa propter mul- 7:47 And Solomon left all the vessels unweighed, because

they were exceeding many: neither was the weight of the brass found out.

titudinem autem nimiam non erat pondus aeris 7:48 fecit-

7:48 And Solomon made all the vessels that pertained unto the house of the LORD: the altar of gold, and the table of gold, whereupon the shewbread was,

que Salomon omnia uasa in domo Domini altare aureum et mensam super quam ponerentur panes propositionis auream 7:49 et candelabra aurea quinque ad dexteram et quinque ad 7:49 And the candlesticks of pure gold, five on the right

side, and five on the left, before the oracle, with the flowers, and the lamps, and the tongs of gold,

sinistram contra oraculum ex auro primo et quasi lilii flores et lucernas desuper aureas et forcipes aureos 7:50 et hyd- 7:50 And the bowls, and the snuffers, and the basons, and

the spoons, and the censers of pure gold; and the hinges of gold, both for the doors of the inner house, the most holy

rias et fuscinulas et fialas et mortariola et turibula de auro place, and for the doors of the house, to wit, of the temple. purissimo et cardines ostiorum domus interioris sancti sanctorum et ostiorum domus templi ex auro erant 7:51 et per- 7:51 So was ended all the work that king Solomon made

for the house of the LORD. And Solomon brought in the things which David his father had dedicated; even the silver, and the gold, and the vessels, did he put among the treasures of the house of the LORD.

fecit omne opus quod faciebat Salomon in domo Domini et intulit quae sanctificauerat Dauid pater suus argentum et aurum et uasa reposuitque in thesauris domus Domini

8:1 tunc congregauit omnes maiores natu Israhel cum prin- 8:1 Then Solomon assembled the elders of Israel, and all

the heads of the tribes, the chief of the fathers of the children of Israel, unto king Solomon in Jerusalem, that they might bring up the ark of the covenant of the LORD out of the city of David, which is Zion.

cipibus tribuum et duces familiarum filiorum Israhel ad regem Salomonem in Hierusalem ut deferrent arcam foederis Domini de ciuitate Dauid id est de Sion 8:2 conuenitque 8:2 And all the men of Israel assembled themselves unto

king Solomon at the feast in the month Ethanim, which is the seventh month.

ad regem Salomonem uniuersus Israhel in mense hethanim in sollemni die ipse est mensis septimus 8:3 ueneruntque 8:3 And all the elders of Israel came, and the priests took

up the ark.

cuncti senes ex Israhel et tulerunt sacerdotes arcam 8:4 et 8:4 And they brought up the ark of the LORD, and the

tabernacle of the congregation, and all the holy vessels that were in the tabernacle, even those did the priests and the

portauerunt arcam Domini et tabernaculum foederis et om- Levites bring up. nia uasa sanctuarii quae erant in tabernaculo et ferebant ea

Biblia Sacra Vulgata

8:5 And king Solomon, and all the congregation of Israel, sacerdotes et Leuitae 8:5 rex autem Salomon et omnis multhat were assembled unto him, were with him before the ark, sacrificing sheep and oxen, that could not be told nor

numbered for multitude. titudo Israhel quae conuenerat ad eum gradiebatur cum illo

ante arcam et immolabant oues et boues absque aestimatione

8:6 And the priests brought in the ark of the covenant of et numero 8:6 et intulerunt sacerdotes arcam foederis Do-

the LORD unto his place, into the oracle of the house, to the most holy place, even under the wings of the cheru-

bims. mini in locum suum in oraculum templi in sanctum sanc-8:7 For the cherubims spread forth their two wings over torum subter alas cherubin 8:7 siquidem cherubin expan-

the place of the ark, and the cherubims covered the ark and the staves thereof above.

debant alas super locum arcae et protegebant arcam et uec-

8:8 And they drew out the staves, that the ends of the tes eius desuper 8:8 cumque eminerent uectes et apparerent

staves were seen out in the holy place before the oracle, and they were not seen without: and there they are unto

this day. summitates eorum foris sanctuarium ante oraculum non ap-

parebant ultra extrinsecus qui et fuerunt ibi usque in prae-

8:9 There was nothing in the ark save the two tables of sentem diem 8:9 in arca autem non est aliud nisi duae ta-

stone, which Moses put there at Horeb, when the LORD made a covenant with the children of Israel, when they

came out of the land of Egypt. bulae lapideae quas posuerat in ea Moses in Horeb quando

pepigit foedus Dominus cum filiis Israhel cum egrederentur

8:10 And it came to pass, when the priests were come out de terra Aegypti 8:10 factum est autem cum exissent sacer-

of the holy place, that the cloud filled the house of the LORD,

dotes de sanctuario nebula impleuit domum Domini 8:11

8:11 So that the priests could not stand to minister because of the cloud: for the glory of the LORD had filled the house of the LORD.

et non poterant sacerdotes stare et ministrare propter nebu-

8:12 Then spake Solomon, The LORD said that he would lam impleuerat enim gloria Domini domum Domini 8:12

dwell in the thick darkness. 8:13 I have surely built thee an house to dwell in, a settled tunc ait Salomon Dominus dixit ut habitaret in nebula 8:13 place for thee to abide in for ever.

aedificans aedificaui domum in habitaculum tuum firmissi-

8:14 And the king turned his face about, and blessed all the mum solium tuum in sempiternum 8:14 conuertitque rex congregation of Israel: (and all the congregation of Israel stood;)

faciem suam et benedixit omni ecclesiae Israhel omnis enim

8:15 And he said, Blessed be the LORD God of Israel, ecclesia Israhel stabat 8:15 et ait benedictus Dominus Deus

which spake with his mouth unto David my father, and hath with his hand fulfilled it, saying,

Israhel qui locutus est ore suo ad Dauid patrem meum et in

8:16 Since the day that I brought forth my people Israel manibus eius perfecit dicens 8:16 a die qua eduxi populum

out of Egypt, I chose no city out of all the tribes of Israel to build an house, that my name might be therein; but I

chose David to be over my people Israel. meum Israhel de Aegypto non elegi ciuitatem de uniuersis tribubus Israhel ut aedificaretur domus et esset nomen meum ibi sed elegi Dauid ut esset super populum meum Israhel

8:17 And it was in the heart of David my father to build an 8:17 uoluitque Dauid pater meus aedificare domum nomini

house for the name of the LORD God of Israel.

8:18 And the LORD said unto David my father, Whereas Domini Dei Israhel 8:18 et ait Dominus ad Dauid patrem

it was in thine heart to build an house unto my name, thou didst well that it was in thine heart.

meum quod cogitasti in corde tuo aedificare domum nomini

8:19 Nevertheless thou shalt not build the house; but thy meo bene fecisti hoc ipsum mente tractans 8:19 uerumtason that shall come forth out of thy loins, he shall build the house unto my name.

men tu non aedificabis domum sed filius tuus qui egredietur

8:20 And the LORD hath performed his word that he de renibus tuis ipse aedificabit domum nomini meo 8:20

spake, and I am risen up in the room of David my father, and sit on the throne of Israel, as the LORD promised, and have built an house for the name of the LORD God of Israel.

confirmauit Dominus sermonem suum quem locutus est stetique pro Dauid patre meo et sedi super thronum Israhel sicut locutus est Dominus et aedificaui domum nomini Domini Dei Israhel 8:21 et constitui ibi locum arcae in qua foedus 8:21 And I have set there a place for the ark, wherein is the

covenant of the LORD, which he made with our fathers, when he brought them out of the land of Egypt.

est Domini quod percussit cum patribus nostris quando egressi sunt de terra Aegypti 8:22 stetit autem Salomon ante 8:22 And Solomon stood before the altar of the LORD in

the presence of all the congregation of Israel, and spread forth his hands toward heaven:

altare Domini in conspectu ecclesiae Israhel et expandit manus suas in caelum 8:23 et ait Domine Deus Israhel non est 8:23 And he said, LORD God of Israel, there is no God

like thee, in heaven above, or on earth beneath, who keepest covenant and mercy with thy servants that walk before

similis tui Deus in caelo desuper et super terra deorsum qui thee with all their heart: custodis pactum et misericordiam seruis tuis qui ambulant coram te in toto corde suo 8:24 qui custodisti seruo tuo Da- 8:24 Who hast kept with thy servant David my father that

thou promisedst him: thou spakest also with thy mouth, and hast fulfilled it with thine hand, as it is this day.

uid patri meo quae locutus es ei ore locutus es et manibus perfecisti ut et haec dies probat 8:25 nunc igitur Domine 8:25 Therefore now, LORD God of Israel, keep with thy

servant David my father that thou promisedst him, saying, There shall not fail thee a man in my sight to sit on the throne of Israel; so that thy children take heed to their way, that they walk before me as thou hast walked before me.

Deus Israhel conserua famulo tuo Dauid patri meo quae locutus es ei dicens non auferetur de te uir coram me qui sedeat super thronum Israhel ita tamen si custodierint filii tui uiam suam ut ambulent coram me sicut tu ambulasti in conspectu meo 8:26 et nunc Deus Israhel firmentur uerba tua 8:26 And now, O God of Israel, let thy word, I pray thee,

be verified, which thou spakest unto thy servant David my father.

quae locutus es seruo tuo Dauid patri meo 8:27 ergone pu-

8:27 But will God indeed dwell on the earth? behold, the heaven and heaven of heavens cannot contain thee; how much less this house that I have builded?

tandum est quod uere Deus habitet super terram si enim caelum et caeli caelorum te capere non possunt quanto magis

domus haec quam aedificaui 8:28 sed respice ad orationem 8:28 Yet have thou respect unto the prayer of thy servant, and to his supplication, O LORD my God, to hearken unto the cry and to the prayer, which thy servant prayeth before

serui tui et ad preces eius Domine Deus meus audi hymnum thee to day: et orationem quam seruus tuus orat coram te hodie 8:29 ut 8:29 That thine eyes may be open toward this house night

and day, even toward the place of which thou hast said, My name shall be there: that thou mayest hearken unto

sint oculi tui aperti super domum hanc nocte et die super the prayer which thy servant shall make toward this place.

domum de qua dixisti erit nomen meum ibi ut exaudias ora-

tionem qua orat te seruus tuus in loco isto 8:30 ut exau- 8:30 And hearken thou to the supplication of thy servant, and of thy people Israel, when they shall pray toward this place: and hear thou in heaven thy dwelling place: and

dias deprecationem serui tui et populi tui Israhel quodcum- when thou hearest, forgive. que orauerint in loco isto et exaudies in loco habitaculi tui in caelo et cum exaudieris propitius eris 8:31 si peccaue- 8:31 If any man trespass against his neighbour, and an oath

be laid upon him to cause him to swear, and the oath come before thine altar in this house:

rit homo in proximum suum et habuerit aliquod iuramentum quo teneatur adstrictus et uenerit propter iuramentum coram altari tuo in domum tuam 8:32 tu exaudies in caelo et fa- 8:32 Then hear thou in heaven, and do, and judge thy ser-

vants, condemning the wicked, to bring his way upon his head; and justifying the righteous, to give him according to his righteousness.

Biblia Sacra Vulgata

cies et iudicabis seruos tuos condemnans impium et reddens uiam suam super caput eius iustificansque iustum et retri-

8:33 When thy people Israel be smitten down before the buens ei secundum iustitiam suam 8:33 si fugerit populus

enemy, because they have sinned against thee, and shall turn again to thee, and confess thy name, and pray, and

make supplication unto thee in this house: tuus Israhel inimicos suos quia peccaturus est tibi et agentes

paenitentiam et confitentes nomini tuo uenerint et orauerint

8:34 Then hear thou in heaven, and forgive the sin of thy et deprecati te fuerint in domo hac 8:34 exaudi in caelo et

people Israel, and bring them again unto the land which thou gavest unto their fathers.

dimitte peccatum populi tui Israhel et reduces eos in terram

8:35 When heaven is shut up, and there is no rain, be- quam dedisti patribus eorum 8:35 si clausum fuerit caelum

cause they have sinned against thee; if they pray toward this place, and confess thy name, and turn from their sin,

when thou afflictest them: et non pluerit propter peccata eorum et orantes in loco isto

paenitentiam egerint nomini tuo et a peccatis suis conuersi

8:36 Then hear thou in heaven, and forgive the sin of thy fuerint propter adflictionem suam 8:36 exaudi eos in caelo

servants, and of thy people Israel, that thou teach them the good way wherein they should walk, and give rain upon thy land, which thou hast given to thy people for an inheritance.

et dimitte peccata seruorum tuorum et populi tui Israhel et ostende eis uiam bonam per quam ambulent et da pluuiam super terram tuam quam dedisti populo tuo in possessionem

8:37 If there be in the land famine, if there be pestilence, 8:37 fames si oborta fuerit in terra aut pestilentia aut corru-

blasting, mildew, locust, or if there be caterpiller; if their enemy besiege them in the land of their cities; whatsoever

plague, whatsoever sickness there be; ptus aer aurugo lucusta rubigo et adflixerit eum et inimicus 8:38 What prayer and supplication soever be made by any eius portas obsidens omnis plaga uniuersa infirmitas 8:38

man, or by all thy people Israel, which shall know every man the plague of his own heart, and spread forth his

hands toward this house: cuncta deuotatio et inprecatio quae acciderit omni homini

de populo tuo Israhel si quis cognouerit plagam cordis sui et

8:39 Then hear thou in heaven thy dwelling place, and expanderit manus suas in domo hac 8:39 tu audies in caelo

forgive, and do, and give to every man according to his ways, whose heart thou knowest; (for thou, even thou only,

knowest the hearts of all the children of men;) in loco habitationis tuae et repropitiaberis et facies ut des

unicuique secundum omnes uias suas sicut uideris cor eius

8:40 That they may fear thee all the days that they live in quia tu nosti solus cor omnium filiorum hominum 8:40 ut the land which thou gavest unto our fathers.

timeant te cunctis diebus quibus uiuunt super faciem terrae

8:41 Moreover concerning a stranger, that is not of thy quam dedisti patribus nostris 8:41 insuper et alienigena qui

people Israel, but cometh out of a far country for thy name’s sake;

non est de populo tuo Israhel cum uenerit de terra longinqua propter nomen tuum audietur enim nomen tuum magnum

8:42 (For they shall hear of thy great name, and of thy et manus tua fortis et brachium tuum 8:42 extentum ubi-

strong hand, and of thy stretched out arm;) when he shall come and pray toward this house;

que cum uenerit ergo et orauerit in loco hoc 8:43 tu exau-

8:43 Hear thou in heaven thy dwelling place, and do according to all that the stranger calleth to thee for: that all people of the earth may know thy name, to fear thee, as do thy people Israel; and that they may know that this house, which I have builded, is called by thy name.

dies in caelo in firmamento habitaculi tui et facies omnia pro quibus inuocauerit te alienigena ut discant uniuersi populi terrarum nomen tuum timere sicut populus tuus Israhel et probent quia nomen tuum inuocatum est super domum hanc quam aedificaui 8:44 si egressus fuerit populus tuus ad bel- 8:44 If thy people go out to battle against their enemy,

whithersoever thou shalt send them, and shall pray unto the LORD toward the city which thou hast chosen, and

lum contra inimicos suos per uiam quocumque miseris eos toward the house that I have built for thy name: orabunt te contra uiam ciuitatis quam elegisti et contra do-

mum quam aedificaui nomini tuo 8:45 et exaudies in caelo 8:45 Then hear thou in heaven their prayer and their sup-

plication, and maintain their cause.

orationem eorum et preces eorum et facies iudicium eorum 8:46 quod si peccauerint tibi non est enim homo qui non pec- 8:46 If they sin against thee, (for there is no man that sin-

neth not,) and thou be angry with them, and deliver them to the enemy, so that they carry them away captives unto

cet et iratus tradideris eos inimicis suis et capti ducti fuerint the land of the enemy, far or near; in terram inimicorum longe uel prope 8:47 et egerint paeni- 8:47 Yet if they shall bethink themselves in the land

whither they were carried captives, and repent, and make supplication unto thee in the land of them that carried them captives, saying, We have sinned, and have done perversely, we have committed wickedness;

tentiam in corde suo in loco captiuitatis et conuersi deprecati te fuerint in captiuitate sua dicentes peccauimus inique egimus impie gessimus 8:48 et reuersi fuerint ad te in uniuerso 8:48 And so return unto thee with all their heart, and with

all their soul, in the land of their enemies, which led them away captive, and pray unto thee toward their land, which thou gavest unto their fathers, the city which thou hast chosen, and the house which I have built for thy name:

corde suo et tota anima sua in terra inimicorum suorum ad quam captiui ducti sunt et orauerint te contra uiam terrae suae quam dedisti patribus eorum et ciuitatis quam elegisti et templi quod aedificaui nomini tuo 8:49 exaudies in caelo 8:49 Then hear thou their prayer and their supplication in

heaven thy dwelling place, and maintain their cause,

in firmamento solii tui orationem eorum et preces et facies iudicium eorum 8:50 et propitiaberis populo tuo qui pecca- 8:50 And forgive thy people that have sinned against thee,

and all their transgressions wherein they have transgressed against thee, and give them compassion before them who carried them captive, that they may have compassion on them:

uit tibi et omnibus iniquitatibus eorum quibus praeuaricati sunt in te et dabis misericordiam coram eis qui eos captiuos habuerint ut misereantur eis 8:51 populus enim tuus est et 8:51 For they be thy people, and thine inheritance, which

thou broughtest forth out of Egypt, from the midst of the furnace of iron:

hereditas tua quos eduxisti de terra Aegypti de medio fornacis ferreae 8:52 ut sint oculi tui aperti ad deprecationem 8:52 That thine eyes may be open unto the supplication of

thy servant, and unto the supplication of thy people Israel, to hearken unto them in all that they call for unto thee.

serui tui et populi tui Israhel et exaudias eos in uniuersis pro quibus inuocauerint te 8:53 tu enim separasti eos tibi in he- 8:53 For thou didst separate them from among all the peo-

ple of the earth, to be thine inheritance, as thou spakest by the hand of Moses thy servant, when thou broughtest our

reditatem de uniuersis populis terrae sicut locutus es per Mo- fathers out of Egypt, O LORD God. sen seruum tuum quando eduxisti patres nostros de Aegypto Domine Deus 8:54 factum est autem cum conplesset Sa- 8:54 And it was so, that when Solomon had made an end

of praying all this prayer and supplication unto the LORD, he arose from before the altar of the LORD, from kneeling

lomon orans Dominum omnem orationem et deprecationem on his knees with his hands spread up to heaven. hanc surrexit de conspectu altaris Domini utrumque enim genu in terram fixerat et manus expanderat ad caelum 8:55 8:55 And he stood, and blessed all the congregation of

Israel with a loud voice, saying,

stetit ergo et benedixit omni ecclesiae Israhel uoce magna dicens 8:56 benedictus Dominus qui dedit requiem populo 8:56 Blessed be the LORD, that hath given rest unto his

people Israel, according to all that he promised: there hath not failed one word of all his good promise, which he

suo Israhel iuxta omnia quae locutus est non cecidit ne unus promised by the hand of Moses his servant.

Biblia Sacra Vulgata

quidem sermo ex omnibus bonis quae locutus est per Mosen

8:57 The LORD our God be with us, as he was with our seruum suum 8:57 sit Dominus Deus noster nobiscum sicut

fathers: let him not leave us, nor forsake us:

fuit cum patribus nostris non derelinquens nos neque proi-

8:58 That he may incline our hearts unto him, to walk in all ciens 8:58 sed inclinet corda nostra ad se ut ambulemus in his ways, and to keep his commandments, and his statutes, and his judgments, which he commanded our fathers.

uniuersis uiis eius et custodiamus mandata eius et caerimo-

8:59 And let these my words, wherewith I have made sup- nias et iudicia quaecumque mandauit patribus nostris 8:59

plication before the LORD, be nigh unto the LORD our God day and night, that he maintain the cause of his servant, and the cause of his people Israel at all times, as the matter shall require:

et sint sermones mei isti quibus deprecatus sum coram Domino adpropinquantes Domino Deo nostro die et nocte ut faciat iudicium seruo suo et populo suo Israhel per singu-

8:60 That all the people of the earth may know that the los dies 8:60 et sciant omnes populi terrae quia Dominus

LORD is God, and that there is none else.

8:61 Let your heart therefore be perfect with the LORD ipse est Deus et non est ultra absque eo 8:61 sit quoque cor

our God, to walk in his statutes, and to keep his commandments, as at this day.

nostrum perfectum cum Domino Deo nostro ut ambulemus in decretis eius et custodiamus mandata eius sicut et hodie

8:62 And the king, and all Israel with him, offered sacrifice 8:62 igitur rex et omnis Israhel cum eo immolabant uictimas

before the LORD.

8:63 And Solomon offered a sacrifice of peace offerings, coram Domino 8:63 mactauitque Salomon hostias pacifi-

which he offered unto the LORD, two and twenty thousand oxen, and an hundred and twenty thousand sheep. So the king and all the children of Israel dedicated the house of the LORD.

cas quas immolauit Domino boum uiginti duo milia ouium centum uiginti milia et dedicauerunt templum Domini rex

8:64 The same day did the king hallow the middle of the et filii Israhel 8:64 in die illa sanctificauit rex medium at-

court that was before the house of the LORD: for there he offered burnt offerings, and meat offerings, and the fat of the peace offerings: because the brasen altar that was before the LORD was too little to receive the burnt offerings, and meat offerings, and the fat of the peace offerings.

rii quod erat ante domum Domini fecit quippe ibi holocaustum et sacrificium et adipem pacificorum quia altare aereum quod erat coram Domino minus erat et capere non poterat

8:65 And at that time Solomon held a feast, and all Is- holocausta et sacrificium et adipem pacificorum 8:65 fecit

rael with him, a great congregation, from the entering in of Hamath unto the river of Egypt, before the LORD our

God, seven days and seven days, even fourteen days. ergo Salomon in tempore illo festiuitatem celebrem et omnis Israhel cum eo multitudo magna ab introitu Emath usque ad riuum Aegypti coram Domino Deo nostro septem diebus

8:66 On the eighth day he sent the people away: and they et septem diebus id est quattuordecim diebus 8:66 et in die

blessed the king, and went unto their tents joyful and glad of heart for all the goodness that the LORD had done for

David his servant, and for Israel his people. octaua dimisit populos qui benedicentes regi profecti sunt in tabernacula sua laetantes et alacri corde super omnibus bonis quae fecerat Dominus Dauid seruo suo et Israhel populo suo

9:1 And it came to pass, when Solomon had finished the 9:1 factum est autem cum perfecisset Salomon aedificium

building of the house of the LORD, and the king’s house, and all Solomon’s desire which he was pleased to do,

domus Domini et aedificium regis et omne quod optauerat et

9:2 That the LORD appeared to Solomon the second time, uoluerat facere 9:2 apparuit Dominus ei secundo sicut ap-

as he had appeared unto him at Gibeon.

paruerat ei in Gabaon 9:3 dixitque Dominus ad eum exau- 9:3 And the LORD said unto him, I have heard thy prayer

and thy supplication, that thou hast made before me: I have hallowed this house, which thou hast built, to put my name there for ever; and mine eyes and mine heart shall be there perpetually.

diui orationem tuam et deprecationem tuam qua deprecatus es coram me sanctificaui domum hanc quam aedificasti ut ponerem nomen meum ibi in sempiternum et erunt oculi mei et cor meum ibi cunctis diebus 9:4 tu quoque si ambulaue- 9:4 And if thou wilt walk before me, as David thy father

walked, in integrity of heart, and in uprightness, to do according to all that I have commanded thee, and wilt keep

ris coram me sicut ambulauit pater tuus in simplicitate cordis my statutes and my judgments: et in aequitate et feceris omnia quae praecepi tibi et legitima mea et iudicia mea seruaueris 9:5 ponam thronum regni tui 9:5 Then I will establish the throne of thy kingdom upon

Israel for ever, as I promised to David thy father, saying, There shall not fail thee a man upon the throne of Israel.

super Israhel in sempiternum sicut locutus sum Dauid patri tuo dicens non auferetur de genere tuo uir de solio Israhel 9:6 si autem auersione auersi fueritis uos et filii uestri non 9:6 But if ye shall at all turn from following me, ye or

your children, and will not keep my commandments and my statutes which I have set before you, but go and serve

sequentes me nec custodientes mandata mea et caerimonias other gods, and worship them: quas proposui uobis sed abieritis et colueritis deos alienos et adoraueritis eos 9:7 auferam Israhel de superficie terrae 9:7 Then will I cut off Israel out of the land which I have

given them; and this house, which I have hallowed for my name, will I cast out of my sight; and Israel shall be a

quam dedi eis et templum quod sanctificaui nomini meo pro- proverb and a byword among all people: iciam a conspectu meo eritque Israhel in prouerbium et in

fabulam cunctis populis 9:8 et domus haec erit in exemp- 9:8 And at this house, which is high, every one that passeth by it shall be astonished, and shall hiss; and they shall say, Why hath the LORD done thus unto this land, and to this

lum omnis qui transierit per eam stupebit et sibilabit et di- house? cet quare fecit Dominus sic terrae huic et domui huic 9:9 9:9 And they shall answer, Because they forsook the

LORD their God, who brought forth their fathers out of the land of Egypt, and have taken hold upon other gods, and have worshipped them, and served them: therefore hath the LORD brought upon them all this evil.

et respondebunt quia dereliquerunt Dominum Deum suum qui eduxit patres eorum de terra Aegypti et secuti sunt deos alienos et adorauerunt eos et coluerunt idcirco induxit Dominus super eos omne malum hoc 9:10 expletis autem an- 9:10 And it came to pass at the end of twenty years, when

Solomon had built the two houses, the house of the LORD, and the king’s house,

nis uiginti postquam aedificauerat Salomon duas domos id

est domum Domini et domum regis 9:11 Hiram rege Tyri 9:11 (Now Hiram the king of Tyre had furnished Solomon with cedar trees and fir trees, and with gold, according to all his desire,) that then king Solomon gave Hiram twenty

praebente Salomoni ligna cedrina et abiegna et aurum iuxta cities in the land of Galilee. omne quod opus habuerat tunc dedit Salomon Hiram uiginti oppida in terra Galileae 9:12 egressusque est Hiram de Tyro 9:12 And Hiram came out from Tyre to see the cities which

Solomon had given him; and they pleased him not.

ut uideret oppida quae dederat ei Salomon et non placuerunt ei 9:13 et ait haecine sunt ciuitates quas dedisti mihi 9:13 And he said, What cities are these which thou hast

given me, my brother? And he called them the land of Cabul unto this day.

frater et appellauit eas terram Chabul usque in diem hanc 9:14 misit quoque Hiram ad regem centum uiginti talenta auri 9:14 And Hiram sent to the king sixscore talents of gold. 9:15 haec est summa expensarum quam obtulit rex Salomon 9:15 And this is the reason of the levy which king Solomon

raised; for to build the house of the LORD, and his own house, and Millo, and the wall of Jerusalem, and Hazor, and Megiddo, and Gezer.

Biblia Sacra Vulgata

ad aedificandam domum Domini et domum suam et Mello et

9:16 For Pharaoh king of Egypt had gone up, and taken murum Hierusalem et Eser et Mageddo et Gazer 9:16 Pha-

Gezer, and burnt it with fire, and slain the Canaanites that dwelt in the city, and given it for a present unto his daugh-

ter, Solomon’s wife. rao rex Aegypti ascendit et cepit Gazer succenditque eam

igni et Chananeum qui habitabat in ciuitate interfecit et de-

9:17 And Solomon built Gezer, and Bethhoron the nether, dit eam in dote filiae suae uxori Salomonis 9:17 aedificauit 9:18 And Baalath, and Tadmor in the wilderness, in the ergo Salomon Gazer et Bethoron inferiorem 9:18 et Baa-

land, 9:19 And all the cities of store that Solomon had, and cities lath et Palmyram in terra solitudinis 9:19 et omnes uicos for his chariots, and cities for his horsemen, and that which Solomon desired to build in Jerusalem, and in Lebanon,

and in all the land of his dominion. qui ad se pertinebant et erant absque muro muniuit et ciuitates curruum et ciuitates equitum et quodcumque ei placuit ut aedificaret in Hierusalem et in Libano et in omni terra

9:20 And all the people that were left of the Amorites, potestatis suae 9:20 uniuersum populum qui remanserat de

Hittites, Perizzites, Hivites, and Jebusites, which were not of the children of Israel,

Amorreis et Hettheis et Ferezeis et Eueis et Iebuseis qui non

9:21 Their children that were left after them in the land, sunt de filiis Israhel 9:21 horum filios qui remanserant in whom the children of Israel also were not able utterly to destroy, upon those did Solomon levy a tribute of bondser-

vice unto this day. terra quos scilicet non potuerant filii Israhel exterminare fe-9:22 But of the children of Israel did Solomon make no cit Salomon tributarios usque ad diem hanc 9:22 de filiis

bondmen: but they were men of war, and his servants, and his princes, and his captains, and rulers of his chariots, and

his horsemen. autem Israhel non constituit Salomon seruire quemquam sed

erant uiri bellatores et ministri eius et principes et duces et

9:23 These were the chief of the officers that were over praefecti curruum et equorum 9:23 erant autem principes

Solomon’s work, five hundred and fifty, which bare rule over the people that wrought in the work.

super omnia opera Salomonis praepositi quingenti quinquaginta qui habebant subiectum populum et statutis operibus

9:24 But Pharaoh’s daughter came up out of the city of imperabant 9:24 filia autem Pharaonis ascendit de ciuitate

David unto her house which Solomon had built for her: then did he build Millo.

Dauid in domum suam quam aedificauerat ei tunc aedifica-

9:25 And three times in a year did Solomon offer burnt uit Mello 9:25 offerebat quoque Salomon tribus uicibus per

offerings and peace offerings upon the altar which he built unto the LORD, and he burnt incense upon the altar that

was before the LORD. So he finished the house. annos singulos holocausta et pacificas uictimas super altare

quod aedificauerat Domino et adolebat thymiama coram Do-

9:26 And king Solomon made a navy of ships in Ezionge- mino perfectumque est templum 9:26 classem quoque fecit

ber, which is beside Eloth, on the shore of the Red sea, in the land of Edom.

rex Salomon in Asiongaber quae est iuxta Ahilam in litore

9:27 And Hiram sent in the navy his servants, shipmen that maris Rubri in terra Idumea 9:27 misitque Hiram in classe had knowledge of the sea, with the servants of Solomon.

illa seruos suos uiros nauticos et gnaros maris cum seruis

9:28 And they came to Ophir, and fetched from thence Salomonis 9:28 qui cum uenissent in Ophir sumptum in-

gold, four hundred and twenty talents, and brought it to king Solomon.

de aurum quadringentorum uiginti talentorum detulerunt ad regem Salomonem

10:1 And when the queen of Sheba heard of the fame of 10:1 sed et regina Saba audita fama Salomonis in nomine

Solomon concerning the name of the LORD, she came to prove him with hard questions.

Domini uenit temptare eum in enigmatibus 10:2 et ingressa 10:2 And she came to Jerusalem with a very great train,

with camels that bare spices, and very much gold, and precious stones: and when she was come to Solomon, she

Hierusalem multo comitatu et diuitiis camelis portantibus communed with him of all that was in her heart. aromata et aurum infinitum nimis et gemmas pretiosas uenit ad Salomonem et locuta est ei uniuersa quae habebat in

corde suo 10:3 et docuit eam Salomon omnia uerba quae 10:3 And Solomon told her all her questions: there was

not any thing hid from the king, which he told her not.

proposuerat non fuit sermo qui regem posset latere et non responderet ei 10:4 uidens autem regina Saba omnem sa- 10:4 And when the queen of Sheba had seen all Solomon’s

wisdom, and the house that he had built,

pientiam Salomonis et domum quam aedificauerat 10:5 et 10:5 And the meat of his table, and the sitting of his servants, and the attendance of his ministers, and their apparel, and his cupbearers, and his ascent by which he went up unto the house of the LORD; there was no more spirit in her.

cibos mensae eius et habitacula seruorum et ordinem ministrantium uestesque eorum et pincernas et holocausta quae offerebat in domo Domini non habebat ultra spiritum 10:6 10:6 And she said to the king, It was a true report that I

heard in mine own land of thy acts and of thy wisdom.

dixitque ad regem uerus est sermo quem audiui in terra mea 10:7 super sermonibus tuis et super sapientia tua et non cre- 10:7 Howbeit I believed not the words, until I came, and

mine eyes had seen it: and, behold, the half was not told me: thy wisdom and prosperity exceedeth the fame which

debam narrantibus mihi donec ipsa ueni et uidi oculis meis I heard. et probaui quod media pars mihi nuntiata non fuerit maior est sapientia et opera tua quam rumor quem audiui 10:8 be- 10:8 Happy are thy men, happy are these thy servants,

which stand continually before thee, and that hear thy wisdom.

ati uiri tui et beati serui tui hii qui stant coram te semper et audiunt sapientiam tuam 10:9 sit Dominus Deus tuus be- 10:9 Blessed be the LORD thy God, which delighted in

thee, to set thee on the throne of Israel: because the LORD loved Israel for ever, therefore made he thee king, to do

nedictus cui placuisti et posuit te super thronum Israhel eo judgment and justice. quod dilexerit Dominus Israhel in sempiternum et constituit te regem ut faceres iudicium et iustitiam 10:10 dedit ergo 10:10 And she gave the king an hundred and twenty tal-

ents of gold, and of spices very great store, and precious stones: there came no more such abundance of spices as

regi centum uiginti talenta auri et aromata multa nimis et these which the queen of Sheba gave to king Solomon. gemmas pretiosas non sunt adlata ultra aromata tam multa

quam ea quae dedit regina Saba regi Salomoni 10:11 sed 10:11 And the navy also of Hiram, that brought gold from Ophir, brought in from Ophir great plenty of almug trees, and precious stones.

et classis Hiram quae portabat aurum de Ophir adtulit ex Ophir ligna thyina multa nimis et gemmas pretiosas 10:12 10:12 And the king made of the almug trees pillars for the

house of the LORD, and for the king’s house, harps also and psalteries for singers: there came no such almug trees,

fecitque rex de lignis thyinis fulchra domus Domini et do- nor were seen unto this day. mus regiae et citharas lyrasque cantoribus non sunt adlata huiuscemodi ligna thyina neque uisa usque in praesentem diem 10:13 rex autem Salomon dedit reginae Saba omnia 10:13 And king Solomon gave unto the queen of Sheba

all her desire, whatsoever she asked, beside that which Solomon gave her of his royal bounty. So she turned and

quae uoluit et petiuit ab eo exceptis his quae ultro obtulerat went to her own country, she and her servants. ei munere regio quae reuersa est et abiit in terram suam cum seruis suis 10:14 erat autem pondus auri quod adferebatur 10:14 Now the weight of gold that came to Solomon in one

year was six hundred threescore and six talents of gold,

Biblia Sacra Vulgata

Salomoni per annos singulos sescentorum sexaginta sex ta-

10:15 Beside that he had of the merchantmen, and of the lentorum auri 10:15 excepto eo quod offerebant uiri qui su-

traffick of the spice merchants, and of all the kings of Arabia, and of the governors of the country.

per uectigalia erant et negotiatores uniuersique scruta uen-

10:16 And king Solomon made two hundred targets of dentes et omnes reges Arabiae ducesque terrae 10:16 fecit

beaten gold: six hundred shekels of gold went to one target.

quoque rex Salomon ducenta scuta de auro puro sescentos

10:17 And he made three hundred shields of beaten gold; auri siclos dedit in lamminas scuti unius 10:17 et trecen-

three pound of gold went to one shield: and the king put them in the house of the forest of Lebanon.

tas peltas ex auro probato trecentae minae auri unam peltam

10:18 Moreover the king made a great throne of ivory, and uestiebant posuitque ea rex in domo siluae Libani 10:18 fe-

overlaid it with the best gold.

cit etiam rex Salomon thronum de ebore grandem et uestiuit

10:19 The throne had six steps, and the top of the throne eum auro fuluo nimis 10:19 qui habebat sex gradus et sum-

was round behind: and there were stays on either side on the place of the seat, and two lions stood beside the stays.

mitas throni rotunda erat in parte posteriori et duae manus hinc atque inde tenentes sedile et duo leones stabant iuxta

10:20 And twelve lions stood there on the one side and on manus singulas 10:20 et duodecim leunculi stantes super

the other upon the six steps: there was not the like made in any kingdom.

sex gradus hinc atque inde non est factum tale opus in uni-

10:21 And all king Solomon’s drinking vessels were of uersis regnis 10:21 sed et omnia uasa de quibus potabat rex

gold, and all the vessels of the house of the forest of Lebanon were of pure gold; none were of silver: it was

nothing accounted of in the days of Solomon. Salomon erant aurea et uniuersa supellex domus saltus Li-

bani de auro purissimo non erat argentum nec alicuius pretii

10:22 For the king had at sea a navy of Tharshish with putabatur in diebus Salomonis 10:22 quia classis regis per

the navy of Hiram: once in three years came the navy of Tharshish, bringing gold, and silver, ivory, and apes, and

peacocks. mare cum classe Hiram semel per tres annos ibat in Thar-

sis deferens inde aurum et argentum dentes elefantorum et

10:23 So king Solomon exceeded all the kings of the earth simias et pauos 10:23 magnificatus est ergo rex Salomon for riches and for wisdom.

10:24 And all the earth sought to Solomon, to hear his super omnes reges terrae diuitiis et sapientia 10:24 et uni-

wisdom, which God had put in his heart.

uersa terra desiderabat uultum Salomonis ut audiret sapien-

10:25 And they brought every man his present, vessels of tiam eius quam dederat Deus in corde eius 10:25 et singuli

silver, and vessels of gold, and garments, and armour, and spices, horses, and mules, a rate year by year.

deferebant ei munera uasa argentea et aurea uestes et arma bellica aromata quoque et equos et mulos per annos singulos

10:26 And Solomon gathered together chariots and horse- 10:26 congregauitque Salomon currus et equites et facti sunt

men: and he had a thousand and four hundred chariots, and twelve thousand horsemen, whom he bestowed in the

cities for chariots, and with the king at Jerusalem. ei mille quadringenti currus et duodecim milia equitum et

disposuit eos per ciuitates munitas et cum rege in Hierusa-

10:27 And the king made silver to be in Jerusalem as lem 10:27 fecitque ut tanta esset abundantia argenti in Hie-

stones, and cedars made he to be as the sycomore trees that are in the vale, for abundance.

rusalem quanta lapidum et cedrorum praebuit multitudinem

10:28 And Solomon had horses brought out of Egypt, and quasi sycomoros quae nascuntur in campestribus 10:28 et

linen yarn: the king’s merchants received the linen yarn at a price.

educebantur equi Salomoni de Aegypto et de Coa negotiatores enim regis emebant de Coa et statuto pretio perducebant 10:29 egrediebatur autem quadriga ex Aegypto sescentis sic- 10:29 And a chariot came up and went out of Egypt for

six hundred shekels of silver, and an horse for an hundred and fifty: and so for all the kings of the Hittites, and for

lis argenti et equus centum quinquaginta atque in hunc mo- the kings of Syria, did they bring them out by their means. dum cuncti reges Hettheorum et Syriae equos uenundabant

11:1 rex autem Salomon amauit mulieres alienigenas mul- 11:1 But king Solomon loved many strange women,

together with the daughter of Pharaoh, women of the Moabites, Ammonites, Edomites, Zidonians, and Hittites:

tas filiam quoque Pharaonis et Moabitidas et Ammanitidas Idumeas et Sidonias et Chettheas 11:2 de gentibus super 11:2 Of the nations concerning which the LORD said unto

the children of Israel, Ye shall not go in to them, neither shall they come in unto you: for surely they will turn away your heart after their gods: Solomon clave unto these in love.

quibus dixit Dominus filiis Israhel non ingrediemini ad eas neque de illis ingredientur ad uestras certissimo enim auertent corda uestra ut sequamini deos earum his itaque copulatus est Salomon ardentissimo amore 11:3 fueruntque ei 11:3 And he had seven hundred wives, princesses, and

three hundred concubines: and his wives turned away his heart.

uxores quasi reginae septingentae et concubinae trecentae et auerterunt mulieres cor eius 11:4 cumque iam esset se- 11:4 For it came to pass, when Solomon was old, that his

wives turned away his heart after other gods: and his heart was not perfect with the LORD his God, as was the heart

nex deprauatum est per mulieres cor eius ut sequeretur deos of David his father. alienos nec erat cor eius perfectum cum Domino Deo suo sicut cor Dauid patris eius 11:5 sed colebat Salomon Ast- 11:5 For Solomon went after Ashtoreth the goddess of the

Zidonians, and after Milcom the abomination of the Ammonites.

harthen deam Sidoniorum et Moloch idolum Ammanitarum 11:6 fecitque Salomon quod non placuerat coram Domino 11:6 And Solomon did evil in the sight of the LORD, and

went not fully after the LORD, as did David his father.

et non adimpleuit ut sequeretur Dominum sicut pater eius 11:7 tunc aedificauit Salomon fanum Chamos idolo Moab in 11:7 Then did Solomon build an high place for Chemosh,

the abomination of Moab, in the hill that is before Jerusalem, and for Molech, the abomination of the chil-

monte qui est contra Hierusalem et Moloch idolo filiorum dren of Ammon. Ammon 11:8 atque in hunc modum fecit uniuersis uxoribus 11:8 And likewise did he for all his strange wives, which

burnt incense and sacrificed unto their gods.

suis alienigenis quae adolebant tura et immolabant diis suis 11:9 igitur iratus est Dominus Salomoni quod auersa esset 11:9 And the LORD was angry with Solomon, because his

heart was turned from the LORD God of Israel, which had appeared unto him twice,

mens eius a Domino Deo Israhel qui apparuerat ei secundo 11:10 et praeceperat de uerbo hoc ne sequeretur deos alienos 11:10 And had commanded him concerning this thing, that

he should not go after other gods: but he kept not that which the LORD commanded.

et non custodiuit quae mandauit ei Dominus 11:11 dixit ita-

11:11 Wherefore the LORD said unto Solomon, Forasmuch as this is done of thee, and thou hast not kept my covenant and my statutes, which I have commanded thee, I will surely rend the kingdom from thee, and will give it to thy servant.

que Dominus Salomoni quia habuisti hoc apud te et non custodisti pactum meum et praecepta mea quae mandaui tibi disrumpens scindam regnum tuum et dabo illud seruo tuo 11:12 uerumtamen in diebus tuis non faciam propter Dauid 11:12 Notwithstanding in thy days I will not do it for

David thy father’s sake: but I will rend it out of the hand of thy son.

patrem tuum de manu filii tui scindam illud 11:13 nec to-

11:13 Howbeit I will not rend away all the kingdom; but will give one tribe to thy son for David my servant’s sake, and for Jerusalem’s sake which I have chosen.

tum regnum auferam sed tribum unam dabo filio tuo pro-

Biblia Sacra Vulgata

11:14 And the LORD stirred up an adversary unto pter Dauid seruum meum et Hierusalem quam elegi 11:14

Solomon, Hadad the Edomite: he was of the king’s seed in Edom.

suscitauit autem Dominus aduersarium Salomoni Adad Idu-

11:15 For it came to pass, when David was in Edom, and meum de semine regio qui erat in Edom 11:15 cum enim Joab the captain of the host was gone up to bury the slain, after he had smitten every male in Edom;

esset Dauid in Idumea et ascendisset Ioab princeps militiae ad sepeliendos eos qui fuerant interfecti et occidisset omne

11:16 (For six months did Joab remain there with all Is- masculinum in Idumea 11:16 sex enim mensibus ibi morarael, until he had cut off every male in Edom:)

tus est Ioab et omnis Israhel donec interimerent omne mas-

11:17 That Hadad fled, he and certain Edomites of his fa- culinum in Idumea 11:17 fugit Adad ipse et uiri idumei de

ther’s servants with him, to go into Egypt; Hadad being yet a little child.

seruis patris eius cum eo ut ingrederetur Aegyptum erat au-

11:18 And they arose out of Midian, and came to Paran: tem Adad puer paruulus 11:18 cumque surrexissent de Ma-

and they took men with them out of Paran, and they came to Egypt, unto Pharaoh king of Egypt; which gave him an

house, and appointed him victuals, and gave him land. dian uenerunt in Pharan tuleruntque secum uiros de Pharan

et introierunt Aegyptum ad Pharaonem regem Aegypti qui

11:19 And Hadad found great favour in the sight of dedit ei domum et cibos constituit et terram delegauit 11:19

Pharaoh, so that he gave him to wife the sister of his own wife, the sister of Tahpenes the queen.

et inuenit Adad gratiam coram Pharao ualde in tantum ut daret ei uxorem sororem uxoris suae germanam Tafnes reginae

11:20 And the sister of Tahpenes bare him Genubath his 11:20 genuitque ei soror Tafnes Genebath filium et nutriuit

son, whom Tahpenes weaned in Pharaoh’s house: and Genubath was in Pharaoh’s household among the sons of

Pharaoh. eum Tafnes in domo Pharaonis eratque Genebath habitans 11:21 And when Hadad heard in Egypt that David slept apud Pharaonem cum filiis eius 11:21 cumque audisset Ad-

with his fathers, and that Joab the captain of the host was dead, Hadad said to Pharaoh, Let me depart, that I may go

to mine own country. ad in Aegypto dormisse Dauid cum patribus suis et mortuum

esse Ioab principem militiae dixit Pharaoni dimitte me ut ua-

11:22 Then Pharaoh said unto him, But what hast thou dam in terram meam 11:22 dixitque ei Pharao qua enim re

lacked with me, that, behold, thou seekest to go to thine own country? And he answered, Nothing: howbeit let me

go in any wise. apud me indiges ut quaeras ire ad terram tuam at ille respon-11:23 And God stirred him up another adversary, Rezon dit nulla sed obsecro ut dimittas me 11:23 suscitauit quo-

the son of Eliadah, which fled from his lord Hadadezer king of Zobah:

que ei Deus aduersarium Razon filium Heliada qui fugerat

11:24 And he gathered men unto him, and became cap- Adadezer regem Soba dominum suum 11:24 et congregauit

tain over a band, when David slew them of Zobah: and they went to Damascus, and dwelt therein, and reigned in

Damascus. contra eum uiros et factus est princeps latronum cum inter-

ficeret eos Dauid abieruntque Damascum et habitauerunt ibi

11:25 And he was an adversary to Israel all the days of et constituerunt eum regem in Damasco 11:25 eratque ad-

Solomon, beside the mischief that Hadad did: and he abhorred Israel, and reigned over Syria.

uersarius Israhel cunctis diebus Salomonis et hoc est malum

11:26 And Jeroboam the son of Nebat, an Ephrathite of Adad et odium contra Israhel regnauitque in Syria 11:26

Zereda, Solomon’s servant, whose mother’s name was Zeruah, a widow woman, even he lifted up his hand

against the king. Hieroboam quoque filius Nabath Ephratheus de Sareda cuius

mater erat nomine Sarua mulier uidua seruus Salomonis le-

11:27 And this was the cause that he lifted up his hand uauit manum contra regem 11:27 et haec causa rebellionis

against the king: Solomon built Millo, and repaired the breaches of the city of David his father.

aduersus eum quia Salomon aedificauit Mello et coaequauit uoraginem ciuitatis Dauid patris sui 11:28 erat autem Hie- 11:28 And the man Jeroboam was a mighty man of valour:

and Solomon seeing the young man that he was industrious, he made him ruler over all the charge of the house of

roboam uir fortis et potens uidensque Salomon adulescen- Joseph. tem bonae indolis et industrium constituerat eum praefectum super tributa uniuersae domus Ioseph 11:29 factum est 11:29 And it came to pass at that time when Jeroboam

went out of Jerusalem, that the prophet Ahijah the Shilonite found him in the way; and he had clad himself

igitur in tempore illo ut Hieroboam egrederetur de Hierusa- with a new garment; and they two were alone in the field: lem et inueniret eum Ahias Silonites propheta in uia opertus pallio nouo erant autem duo tantum in agro 11:30 ad- 11:30 And Ahijah caught the new garment that was on

him, and rent it in twelve pieces:

prehendensque Ahia pallium suum nouum quo opertus erat scidit in duodecim partes 11:31 et ait ad Hieroboam tolle 11:31 And he said to Jeroboam, Take thee ten pieces: for

thus saith the LORD, the God of Israel, Behold, I will rend the kingdom out of the hand of Solomon, and will give ten

tibi decem scissuras haec enim dicit Dominus Deus Israhel tribes to thee: ecce ego scindam regnum de manu Salomonis et dabo tibi decem tribus 11:32 porro una tribus remanebit ei propter 11:32 (But he shall have one tribe for my servant David’s

sake, and for Jerusalem’s sake, the city which I have chosen out of all the tribes of Israel:)

seruum meum Dauid et Hierusalem ciuitatem quam elegi ex omnibus tribubus Israhel 11:33 eo quod dereliquerint me et 11:33 Because that they have forsaken me, and have wor-

shipped Ashtoreth the goddess of the Zidonians, Chemosh the god of the Moabites, and Milcom the god of the children of Ammon, and have not walked in my ways, to do that which is right in mine eyes, and to keep my statutes and my judgments, as did David his father.

adorauerint Astharoth deam Sidoniorum et Chamos deum Moab et Melchom deum filiorum Ammon et non ambulauerint in uiis meis ut facerent iustitiam coram me et praecepta mea et iudicia sicut Dauid pater eius 11:34 nec au- 11:34 Howbeit I will not take the whole kingdom out of

his hand: but I will make him prince all the days of his life for David my servant’s sake, whom I chose, because

feram omne regnum de manu eius sed ducem ponam eum he kept my commandments and my statutes: cunctis diebus uitae suae propter Dauid seruum meum quem elegi qui custodiuit mandata mea et praecepta mea 11:35 11:35 But I will take the kingdom out of his son’s hand,

and will give it unto thee, even ten tribes.

auferam autem regnum de manu filii eius et dabo tibi de-

cem tribus 11:36 filio autem eius dabo tribum unam ut re- 11:36 And unto his son will I give one tribe, that David my servant may have a light alway before me in Jerusalem, the city which I have chosen me to put my name there.

maneat lucerna Dauid seruo meo cunctis diebus coram me in Hierusalem ciuitatem quam elegi ut esset nomen meum

ibi 11:37 te autem adsumam et regnabis super omnia quae 11:37 And I will take thee, and thou shalt reign according

to all that thy soul desireth, and shalt be king over Israel.

desiderat anima tua erisque rex super Israhel 11:38 si igitur 11:38 And it shall be, if thou wilt hearken unto all that I

command thee, and wilt walk in my ways, and do that is right in my sight, to keep my statutes and my commandments, as David my servant did; that I will be with thee, and build thee a sure house, as I built for David, and will give Israel unto thee.

audieris omnia quae praecepero tibi et ambulaueris in uiis meis et feceris quod rectum est coram me custodiens mandata mea et praecepta mea sicut fecit Dauid seruus meus ero tecum et aedificabo tibi domum fidelem quomodo aedificaui Dauid et tradam tibi Israhel 11:39 et adfligam semen Da- 11:39 And I will for this afflict the seed of David, but not

for ever.

Biblia Sacra Vulgata

11:40 Solomon sought therefore to kill Jeroboam. And uid super hoc uerumtamen non cunctis diebus 11:40 uoluit

Jeroboam arose, and fled into Egypt, unto Shishak king of Egypt, and was in Egypt until the death of Solomon.

ergo Salomon interficere Hieroboam qui surrexit et aufugit in Aegyptum ad Susac regem Aegypti et fuit in Aegypto us-

11:41 And the rest of the acts of Solomon, and all that he que ad mortem Salomonis 11:41 reliquum autem uerborum did, and his wisdom, are they not written in the book of the acts of Solomon?

Salomonis et omnia quae fecit et sapientia eius ecce uni-

11:42 And the time that Solomon reigned in Jerusalem uersa scripta sunt in libro uerborum Salomonis 11:42 dies

over all Israel was forty years.

autem quos regnauit Salomon in Hierusalem super omnem

11:43 And Solomon slept with his fathers, and was buried Israhel quadraginta anni sunt 11:43 dormiuitque Salomon

in the city of David his father: and Rehoboam his son reigned in his stead.

cum patribus suis et sepultus est in ciuitate Dauid patris sui regnauitque Roboam filius eius pro eo

12:1 And Rehoboam went to Shechem: for all Israel were 12:1 uenit autem Roboam in Sychem illuc enim congrega-

come to Shechem to make him king.

12:2 And it came to pass, when Jeroboam the son of tus erat omnis Israhel ad constituendum eum regem 12:2 at

Nebat, who was yet in Egypt, heard of it, (for he was fled from the presence of king Solomon, and Jeroboam dwelt

in Egypt;) Hieroboam filius Nabath cum adhuc esset in Aegypto pro-

fugus a facie regis Salomonis audita morte eius reuersus est

12:3 That they sent and called him. And Jeroboam and de Aegypto 12:3 miseruntque et uocauerunt eum uenit ergo

all the congregation of Israel came, and spake unto Rehoboam, saying,

Hieroboam et omnis multitudo Israhel et locuti sunt ad Ro-

12:4 Thy father made our yoke grievous: now therefore boam dicentes 12:4 pater tuus durissimum iugum inposuit

make thou the grievous service of thy father, and his heavy yoke which he put upon us, lighter, and we will serve thee.

nobis tu itaque nunc inminue paululum de imperio patris tui durissimo et de iugo grauissimo quod inposuit nobis et ser-

12:5 And he said unto them, Depart yet for three days, uiemus tibi 12:5 qui ait eis ite usque ad tertium diem et

then come again to me. And the people departed.

12:6 And king Rehoboam consulted with the old men, that reuertimini ad me cumque abisset populus 12:6 iniit consi-

stood before Solomon his father while he yet lived, and said, How do ye advise that I may answer this people?

lium rex Roboam cum senibus qui adsistebant coram Salomone patre eius dum aduiueret et ait quod mihi datis consi-

12:7 And they spake unto him, saying, If thou wilt be a lium ut respondeam populo 12:7 qui dixerunt ei si hodie ob-

servant unto this people this day, and wilt serve them, and answer them, and speak good words to them, then they

will be thy servants for ever. oedieris populo huic et seruieris et petitioni eorum cesseris

locutusque fueris ad eos uerba lenia erunt tibi serui cunctis

12:8 But he forsook the counsel of the old men, which diebus 12:8 qui dereliquit consilium senum quod dederant

they had given him, and consulted with the young men that were grown up with him, and which stood before him:

ei et adhibuit adulescentes qui nutriti fuerant cum eo et adsis-

12:9 And he said unto them, What counsel give ye that we tebant illi 12:9 dixitque ad eos quod mihi datis consilium ut

may answer this people, who have spoken to me, saying, Make the yoke which thy father did put upon us lighter?

respondeam populo huic qui dixerunt mihi leuius fac iugum

12:10 And the young men that were grown up with him quod inposuit pater tuus super nos 12:10 et dixerunt ei iuue-

spake unto him, saying, Thus shalt thou speak unto this people that spake unto thee, saying, Thy father made our yoke heavy, but make thou it lighter unto us; thus shalt thou say unto them, My little finger shall be thicker than my father’s loins.

nes qui nutriti fuerant cum eo sic loquere populo huic qui locuti sunt ad te dicentes pater tuus adgrauauit iugum nostrum tu releua nos sic loqueris ad eos minimus digitus meus grossior est dorso patris mei 12:11 et nunc pater meus posuit su- 12:11 And now whereas my father did lade you with a

heavy yoke, I will add to your yoke: my father hath chastised you with whips, but I will chastise you with scorpi-

per uos iugum graue ego autem addam super iugum uestrum ons. pater meus cecidit uos flagellis ego autem caedam scorpionibus 12:12 uenit ergo Hieroboam et omnis populus ad Ro- 12:12 So Jeroboam and all the people came to Rehoboam

the third day, as the king had appointed, saying, Come to me again the third day.

boam die tertia sicut locutus fuerat rex dicens reuertimini ad me die tertia 12:13 responditque rex populo dura derelicto 12:13 And the king answered the people roughly, and for-

sook the old men’s counsel that they gave him;

consilio seniorum quod ei dederant 12:14 et locutus est eis 12:14 And spake to them after the counsel of the young

men, saying, My father made your yoke heavy, and I will add to your yoke: my father also chastised you with whips,

secundum consilium iuuenum dicens pater meus adgrauauit but I will chastise you with scorpions. iugum uestrum ego autem addam iugo uestro pater meus cecidit uos flagellis et ego caedam scorpionibus 12:15 et non 12:15 Wherefore the king hearkened not unto the people;

for the cause was from the LORD, that he might perform his saying, which the LORD spake by Ahijah the Shilonite

adquieuit rex populo quoniam auersatus eum fuerat Domi- unto Jeroboam the son of Nebat. nus ut suscitaret uerbum suum quod locutus fuerat in manu Ahiae Silonitae ad Hieroboam filium Nabath 12:16 uidens 12:16 So when all Israel saw that the king hearkened not

unto them, the people answered the king, saying, What portion have we in David? neither have we inheritance in the son of Jesse: to your tents, O Israel: now see to thine own house, David. So Israel departed unto their tents.

itaque populus quod noluisset eos audire rex respondit ei dicens quae nobis pars in Dauid uel quae hereditas in filio Isai in tabernacula tua Israhel nunc uide domum tuam Dauid et abiit Israhel in tabernacula sua 12:17 super filios autem Is- 12:17 But as for the children of Israel which dwelt in the

cities of Judah, Rehoboam reigned over them.

rahel quicumque habitabant in ciuitatibus Iuda regnauit Roboam 12:18 misit igitur rex Roboam Aduram qui erat super 12:18 Then king Rehoboam sent Adoram, who was over

the tribute; and all Israel stoned him with stones, that he died. Therefore king Rehoboam made speed to get him up

tributum et lapidauit eum omnis Israhel et mortuus est porro to his chariot, to flee to Jerusalem. rex Roboam festinus ascendit currum et fugit in Hierusalem 12:19 recessitque Israhel a domo Dauid usque in praesentem 12:19 So Israel rebelled against the house of David unto

this day.

diem 12:20 factum est autem cum audisset omnis Israhel 12:20 And it came to pass, when all Israel heard that Jeroboam was come again, that they sent and called him unto the congregation, and made him king over all Israel: there was none that followed the house of David, but the tribe of Judah only.

quod reuersus esset Hieroboam miserunt et uocauerunt eum congregato coetu et constituerunt regem super omnem Israhel nec secutus est quisquam domum Dauid praeter tribum Iuda solam 12:21 uenit autem Roboam Hierusalem et con- 12:21 And when Rehoboam was come to Jerusalem, he as-

sembled all the house of Judah, with the tribe of Benjamin, an hundred and fourscore thousand chosen men, which were warriors, to fight against the house of Israel, to bring the kingdom again to Rehoboam the son of Solomon.

gregauit uniuersam domum Iuda et tribum Beniamin centum octoginta milia electorum uirorum et bellatorum ut pugnaret contra domum Israhel et reduceret regnum Roboam filio Salomonis 12:22 factus est uero sermo Domini ad Semeiam 12:22 But the word of God came unto Shemaiah the man

of God, saying,

uirum Dei dicens 12:23 loquere ad Roboam filium Salomo- 12:23 Speak unto Rehoboam, the son of Solomon, king of

Judah, and unto all the house of Judah and Benjamin, and to the remnant of the people, saying,

Biblia Sacra Vulgata

nis regem Iuda et ad omnem domum Iuda et Beniamin et

12:24 Thus saith the LORD, Ye shall not go up, nor fight reliquos de populo dicens 12:24 haec dicit Dominus non

against your brethren the children of Israel: return every man to his house; for this thing is from me. They hearkened therefore to the word of the LORD, and returned to depart, according to the word of the LORD.

ascendetis nec bellabitis contra fratres uestros filios Israhel reuertatur uir in domum suam a me enim factum est uerbum hoc audierunt sermonem Domini et reuersi sunt de itinere

12:25 Then Jeroboam built Shechem in mount Ephraim, sicut eis praeceperat Dominus 12:25 aedificauit autem Hie-

and dwelt therein; and went out from thence, and built Penuel.

roboam Sychem in monte Ephraim et habitauit ibi et egres-

12:26 And Jeroboam said in his heart, Now shall the king- sus inde aedificauit Phanuhel 12:26 dixitque Hieroboam in

dom return to the house of David: 12:27 If this people go up to do sacrifice in the house of corde suo nunc reuertetur regnum ad domum Dauid 12:27 the LORD at Jerusalem, then shall the heart of this people turn again unto their lord, even unto Rehoboam king of Judah, and they shall kill me, and go again to Rehoboam king of Judah.

si ascenderit populus iste ut faciat sacrificia in domo Domini in Hierusalem et conuertetur cor populi huius ad dominum suum Roboam regem Iuda interficientque me et reuer-

12:28 Whereupon the king took counsel, and made two tentur ad eum 12:28 et excogitato consilio fecit duos uitu-

calves of gold, and said unto them, It is too much for you to go up to Jerusalem: behold thy gods, O Israel, which

brought thee up out of the land of Egypt. los aureos et dixit eis nolite ultra ascendere Hierusalem ecce 12:29 And he set the one in Bethel, and the other put he in dii tui Israhel qui eduxerunt te de terra Aegypti 12:29 po-

Dan. 12:30 And this thing became a sin: for the people went to suitque unum in Bethel et alterum in Dan 12:30 et factum worship before the one, even unto Dan.

est uerbum hoc in peccatum ibat enim populus ad adoran-

12:31 And he made an house of high places, and made dum uitulum usque in Dan 12:31 et fecit fana in excelsis

priests of the lowest of the people, which were not of the sons of Levi.

et sacerdotes de extremis populi qui non erant de filiis Leui

12:32 And Jeroboam ordained a feast in the eighth month, 12:32 constituitque diem sollemnem in mense octauo quinta-

on the fifteenth day of the month, like unto the feast that is in Judah, and he offered upon the altar. So did he in Bethel, sacrificing unto the calves that he had made: and he placed in Bethel the priests of the high places which he had made.

decima die mensis in similitudinem sollemnitatis quae celebratur in Iuda et ascendens altare similiter fecit in Bethel ut immolaret uitulis quos fabricatus erat constituitque in Bethel

12:33 So he offered upon the altar which he had made sacerdotes excelsorum quae fecerat 12:33 et ascendit super

in Bethel the fifteenth day of the eighth month, even in the month which he had devised of his own heart; and ordained a feast unto the children of Israel: and he offered upon the altar, and burnt incense.

altare quod extruxerat in Bethel quintadecima die mensis octaui quem finxerat de corde suo et fecit sollemnitatem filiis Israhel et ascendit super altare ut adoleret incensum

13:1 And, behold, there came a man of God out of Ju- 13:1 et ecce uir Dei uenit de Iuda in sermone Domini in

dah by the word of the LORD unto Bethel: and Jeroboam stood by the altar to burn incense.

Bethel Hieroboam stante super altare et tus iaciente 13:2 et

13:2 And he cried against the altar in the word of the LORD, and said, O altar, altar, thus saith the LORD; Behold, a child shall be born unto the house of David, Josiah by name; and upon thee shall he offer the priests of the high places that burn incense upon thee, and men’s bones shall be burnt upon thee.

exclamauit contra altare in sermone Domini et ait altare altare haec dicit Dominus ecce filius nascetur domui Dauid Iosias nomine et immolabit super te sacerdotes excelsorum qui nunc in te tura succendunt et ossa hominum incendet super te 13:3 deditque in die illa signum dicens hoc erit signum 13:3 And he gave a sign the same day, saying, This is the

sign which the LORD hath spoken; Behold, the altar shall be rent, and the ashes that are upon it shall be poured out.

quod locutus est Dominus ecce altare scinditur et effunditur cinis qui in eo est 13:4 cumque audisset rex sermonem ho- 13:4 And it came to pass, when king Jeroboam heard the

saying of the man of God, which had cried against the altar in Bethel, that he put forth his hand from the altar, saying, Lay hold on him. And his hand, which he put forth against him, dried up, so that he could not pull it in again to him.

minis Dei quem inclamauerat contra altare in Bethel extendit manum suam de altari dicens adprehendite eum et exaruit manus eius quam extenderat contra eum nec ualuit retrahere

eam ad se 13:5 altare quoque scissum est et effusus cinis 13:5 The altar also was rent, and the ashes poured out from the altar, according to the sign which the man of God had given by the word of the LORD.

de altari iuxta signum quod praedixerat uir Dei in sermone Domini 13:6 et ait rex ad uirum Dei deprecare faciem Do- 13:6 And the king answered and said unto the man of God,

Intreat now the face of the LORD thy God, and pray for me, that my hand may be restored me again. And the man of God besought the LORD, and the king’s hand was restored him again, and became as it was before.

mini Dei tui et ora pro me ut restituatur manus mea mihi orauit uir Dei faciem Domini et reuersa est manus regis ad eum et facta est sicut prius fuerat 13:7 locutus est autem 13:7 And the king said unto the man of God, Come home

with me, and refresh thyself, and I will give thee a reward.

rex ad uirum Dei ueni mecum domum ut prandeas et dabo tibi munera 13:8 responditque uir Dei ad regem si dede- 13:8 And the man of God said unto the king, If thou wilt

give me half thine house, I will not go in with thee, neither will I eat bread nor drink water in this place:

ris mihi mediam partem domus tuae non ueniam tecum nec comedam panem neque bibam aquam in loco isto 13:9 sic 13:9 For so was it charged me by the word of the LORD,

saying, Eat no bread, nor drink water, nor turn again by the same way that thou camest.

enim mandatum est mihi in sermone Domini praecipientis non comedes panem neque bibes aquam nec reuerteris per uiam qua uenisti 13:10 abiit ergo per aliam uiam et non est 13:10 So he went another way, and returned not by the

way that he came to Bethel.

reuersus per iter quo uenerat in Bethel 13:11 prophetes au- 13:11 Now there dwelt an old prophet in Bethel; and his

sons came and told him all the works that the man of God had done that day in Bethel: the words which he had spo-

tem quidam senex habitabat in Bethel ad quem uenit filius ken unto the king, them they told also to their father. suus et narrauit ei omnia opera quae fecerat uir Dei illa die in Bethel et uerba quae locutus fuerat ad regem et narrauerunt patri suo 13:12 et dixit eis pater eorum per quam uiam 13:12 And their father said unto them, What way went

he? For his sons had seen what way the man of God went, which came from Judah.

abiit ostenderunt ei filii sui uiam per quam abierat uir Dei qui uenerat de Iuda 13:13 et ait filiis suis sternite mihi asinum 13:13 And he said unto his sons, Saddle me the ass. So

they saddled him the ass: and he rode thereon,

qui cum strauissent ascendit 13:14 et abiit post uirum Dei 13:14 And went after the man of God, and found him sit-

ting under an oak: and he said unto him, Art thou the man of God that camest from Judah? And he said, I am.

et inuenit eum sedentem subtus terebinthum et ait illi tune es uir Dei qui uenisti de Iuda respondit ille ego sum 13:15 13:15 Then he said unto him, Come home with me, and

eat bread.

dixit ad eum ueni mecum domum ut comedas panem 13:16 13:16 And he said, I may not return with thee, nor go in with thee: neither will I eat bread nor drink water with thee in this place:

qui ait non possum reuerti neque uenire tecum nec comedam panem nec bibam aquam in loco isto 13:17 quia locu- 13:17 For it was said to me by the word of the LORD,

Thou shalt eat no bread nor drink water there, nor turn again to go by the way that thou camest.

tus est Dominus ad me in sermone Domini dicens non co-

Biblia Sacra Vulgata

medes panem et non bibes ibi aquam nec reuerteris per uiam

13:18 He said unto him, I am a prophet also as thou art; qua ieris 13:18 qui ait illi et ego propheta sum similis tui

and an angel spake unto me by the word of the LORD, saying, Bring him back with thee into thine house, that he

may eat bread and drink water. But he lied unto him. et angelus locutus est mihi in sermone Domini dicens re-

duc eum tecum in domum tuam et comedat panem et bibat

13:19 So he went back with him, and did eat bread in his aquam fefellit eum 13:19 et reduxit secum comedit ergo pa-

house, and drank water.

13:20 And it came to pass, as they sat at the table, that nem in domo eius et bibit aquam 13:20 cumque sederent ad

the word of the LORD came unto the prophet that brought him back:

mensam factus est sermo Domini ad prophetam qui reduxe-

13:21 And he cried unto the man of God that came from rat eum 13:21 et exclamauit ad uirum Dei qui uenerat de

Judah, saying, Thus saith the LORD, Forasmuch as thou hast disobeyed the mouth of the LORD, and hast not kept the commandment which the LORD thy God commanded thee,

Iuda dicens haec dicit Dominus quia inoboediens fuisti ori Domini et non custodisti mandatum quod praecepit tibi Do-

13:22 But camest back, and hast eaten bread and drunk minus Deus tuus 13:22 et reuersus es et comedisti panem

water in the place, of the which the Lord did say to thee, Eat no bread, and drink no water; thy carcase shall not

come unto the sepulchre of thy fathers. et bibisti aquam in loco in quo praecepit tibi ne comede-

res panem neque biberes aquam non inferetur cadauer tuum

13:23 And it came to pass, after he had eaten bread, and in sepulchrum patrum tuorum 13:23 cumque comedisset et

after he had drunk, that he saddled for him the ass, to wit, for the prophet whom he had brought back.

bibisset strauit asinum prophetae quem reduxerat 13:24 qui

13:24 And when he was gone, a lion met him by the way, and slew him: and his carcase was cast in the way, and the ass stood by it, the lion also stood by the carcase.

cum abisset inuenit eum leo in uia et occidit et erat cadauer eius proiectum in itinere asinus autem stabat iuxta illum et

13:25 And, behold, men passed by, and saw the carcase leo stabat iuxta cadauer 13:25 et ecce uiri transeuntes uide-

cast in the way, and the lion standing by the carcase: and they came and told it in the city where the old prophet

dwelt. runt cadauer proiectum in uia et leonem stantem iuxta cada-

uer et uenerunt et diuulgauerunt in ciuitate in qua prophetes

13:26 And when the prophet that brought him back from senex ille habitabat 13:26 quod cum audisset propheta ille

the way heard thereof, he said, It is the man of God, who was disobedient unto the word of the LORD: therefore the LORD hath delivered him unto the lion, which hath torn him, and slain him, according to the word of the LORD, which he spake unto him.

qui reduxerat eum de uia ait uir Dei est qui inoboediens fuit ori Domini et tradidit eum Dominus leoni et confregit eum

13:27 And he spake to his sons, saying, Saddle me the ass. et occidit iuxta uerbum Domini quod locutus est ei 13:27 And they saddled him.

dixitque ad filios suos sternite mihi asinum qui cum strauis-

13:28 And he went and found his carcase cast in the way, sent 13:28 et ille abisset inuenit cadauer eius proiectum in

and the ass and the lion standing by the carcase: the lion had not eaten the carcase, nor torn the ass.

uia et asinum et leonem stantes iuxta cadauer non comedit

13:29 And the prophet took up the carcase of the man of leo de cadauere nec laesit asinum 13:29 tulit ergo prophe-

God, and laid it upon the ass, and brought it back: and the old prophet came to the city, to mourn and to bury him.

tes cadauer uiri Dei et posuit illud super asinum et reuersus intulit in ciuitatem prophetae senis ut plangerent eum

13:30 And he laid his carcase in his own grave; and they 13:30 et posuit cadauer eius in sepulchro suo et planxerunt

mourned over him, saying, Alas, my brother! 13:31 And it came to pass, after he had buried him, that he eum heu frater 13:31 cumque planxissent eum dixit ad fispake to his sons, saying, When I am dead, then bury me in the sepulchre wherein the man of God is buried; lay my

bones beside his bones: lios suos cum mortuus fuero sepelite me in sepulchro in quo uir Dei sepultus est iuxta ossa eius ponite ossa mea 13:32 13:32 For the saying which he cried by the word of

the LORD against the altar in Bethel, and against all the houses of the high places which are in the cities of

profecto enim ueniet sermo quem praedixit in sermone Do- Samaria, shall surely come to pass. mini contra altare quod est in Bethel et contra omnia fana excelsorum quae sunt in urbibus Samariae 13:33 post uerba 13:33 After this thing Jeroboam returned not from his evil

way, but made again of the lowest of the people priests of the high places: whosoever would, he consecrated him,

haec non est reuersus Hieroboam de uia sua pessima sed e and he became one of the priests of the high places. contrario fecit de nouissimis populi sacerdotes excelsorum quicumque uolebat implebat manum suam et fiebat sacerdos excelsorum 13:34 et propter hanc causam peccauit domus 13:34 And this thing became sin unto the house of Jer-

oboam, even to cut it off, and to destroy it from off the face of the earth.

Hieroboam et euersa est et deleta de superficie terrae

14:1 in tempore illo aegrotauit Abia filius Hieroboam 14:1 At that time Abijah the son of Jeroboam fell sick.

14:2 dixitque Hieroboam uxori suae surge et commuta ha- 14:2 And Jeroboam said to his wife, Arise, I pray thee, and disguise thyself, that thou be not known to be the wife of Jeroboam; and get thee to Shiloh: behold, there is Ahijah the prophet, which told me that I should be king over this people.

bitum ne cognoscaris quod sis uxor Hieroboam et uade in Silo ubi est Ahia propheta qui locutus est mihi quod regnaturus essem super populum hunc 14:3 tolle quoque in manu 14:3 And take with thee ten loaves, and cracknels, and a

cruse of honey, and go to him: he shall tell thee what shall become of the child.

tua decem panes et crustula et uas mellis et uade ad illum ipse indicabit tibi quid euenturum sit huic puero 14:4 fe- 14:4 And Jeroboam’s wife did so, and arose, and went to

Shiloh, and came to the house of Ahijah. But Ahijah could not see; for his eyes were set by reason of his age.

cit ut dixerat uxor Hieroboam et consurgens abiit in Silo et uenit in domum Ahia at ille non poterat uidere quia caligauerant oculi eius prae senectute 14:5 dixit autem Dominus ad 14:5 And the LORD said unto Ahijah, Behold, the wife of

Jeroboam cometh to ask a thing of thee for her son; for he is sick: thus and thus shalt thou say unto her: for it shall be, when she cometh in, that she shall feign herself to be another woman.

Ahiam ecce uxor Hieroboam ingreditur ut consulat te super filio suo qui aegrotat haec et haec loqueris ei cum ergo illa intraret et dissimularet se esse quae erat 14:6 audiuit Ahias 14:6 And it was so, when Ahijah heard the sound of her

feet, as she came in at the door, that he said, Come in, thou wife of Jeroboam; why feignest thou thyself to be another?

sonitum pedum eius introeuntis per ostium et ait ingredere for I am sent to thee with heavy tidings. uxor Hieroboam quare aliam esse te simulas ego autem missus sum ad te durus nuntius 14:7 uade et dic Hieroboam 14:7 Go, tell Jeroboam, Thus saith the LORD God of Is-

rael, Forasmuch as I exalted thee from among the people, and made thee prince over my people Israel,

haec dicit Dominus Deus Israhel quia exaltaui te de medio populi et dedi te ducem super populum meum Israhel 14:8 14:8 And rent the kingdom away from the house of David,

and gave it thee: and yet thou hast not been as my servant David, who kept my commandments, and who followed me with all his heart, to do that only which was right in mine eyes;

et scidi regnum domus Dauid et dedi illud tibi et non fuisti sicut seruus meus Dauid qui custodiuit mandata mea et secutus est me in toto corde suo faciens quod placitum esset

in conspectu meo 14:9 sed operatus es male super omnes 14:9 But hast done evil above all that were before thee: for thou hast gone and made thee other gods, and molten images, to provoke me to anger, and hast cast me behind

qui fuerunt ante te et fecisti tibi deos alienos et conflatiles ut thy back: me ad iracundiam prouocares me autem proiecisti post cor-

Biblia Sacra Vulgata

14:10 Therefore, behold, I will bring evil upon the house pus tuum 14:10 idcirco ecce ego inducam mala super doof Jeroboam, and will cut off from Jeroboam him that pisseth against the wall, and him that is shut up and left in Israel, and will take away the remnant of the house of Jeroboam, as a man taketh away dung, till it be all gone.

mum Hieroboam et percutiam de Hieroboam mingentem ad parietem et clausum et nouissimum in Israhel et mundabo reliquias domus Hieroboam sicut mundari solet fimus usque

14:11 Him that dieth of Jeroboam in the city shall the dogs ad purum 14:11 qui mortui fuerint de Hieroboam in ciuieat; and him that dieth in the field shall the fowls of the air eat: for the LORD hath spoken it.

tate comedent eos canes qui autem mortui fuerint in agro

14:12 Arise thou therefore, get thee to thine own house: uorabunt eos aues caeli quia Dominus locutus est 14:12 tu

and when thy feet enter into the city, the child shall die.

igitur surge et uade in domum tuam et in ipso introitu pe-

14:13 And all Israel shall mourn for him, and bury him: dum tuorum in urbem morietur puer 14:13 et planget eum for he only of Jeroboam shall come to the grave, because in him there is found some good thing toward the LORD

God of Israel in the house of Jeroboam. omnis Israhel et sepeliet iste enim solus infertur de Hiero-

boam in sepulchrum quia inuentus est super eo sermo bonus

14:14 Moreover the LORD shall raise him up a king over ad Dominum Deum Israhel in domo Hieroboam 14:14 con-Israel, who shall cut off the house of Jeroboam that day: but what? even now.

stituet autem sibi Dominus regem super Israhel qui percu-

14:15 For the LORD shall smite Israel, as a reed is shaken tiat domum Hieroboam in hac die et in hoc tempore 14:15 in the water, and he shall root up Israel out of this good land, which he gave to their fathers, and shall scatter them beyond the river, because they have made their groves, provoking the LORD to anger.

et percutiet Dominus Israhel sicut moueri solet harundo in aqua et euellet Israhel de terra bona hac quam dedit patribus eorum et uentilabit eos trans Flumen quia fecerunt sibi

14:16 And he shall give Israel up because of the sins of lucos ut inritarent Dominum 14:16 et tradet Dominus Isra-

Jeroboam, who did sin, and who made Israel to sin.

hel propter peccata Hieroboam qui peccauit et peccare fe-

14:17 And Jeroboam’s wife arose, and departed, and came cit Israhel 14:17 surrexit itaque uxor Hieroboam et abiit

to Tirzah: and when she came to the threshold of the door, the child died;

et uenit in Thersa cumque illa ingrederetur limen domus

14:18 And they buried him; and all Israel mourned for puer mortuus est 14:18 et sepelierunt eum et planxit illum

him, according to the word of the LORD, which he spake by the hand of his servant Ahijah the prophet.

omnis Israhel iuxta sermonem Domini quem locutus est in

14:19 And the rest of the acts of Jeroboam, how he warred, manu serui sui Ahiae prophetae 14:19 reliqua autem uerbo-

and how he reigned, behold, they are written in the book of the chronicles of the kings of Israel.

rum Hieroboam quomodo pugnauerit et quomodo regnauerit ecce scripta sunt in libro uerborum dierum regum Israhel

14:20 And the days which Jeroboam reigned were two and 14:20 dies autem quibus regnauit Hieroboam uiginti duo anni

twenty years: and he slept with his fathers, and Nadab his son reigned in his stead.

sunt et dormiuit cum patribus suis regnauitque Nadab filius

14:21 And Rehoboam the son of Solomon reigned in Ju- eius pro eo 14:21 porro Roboam filius Salomonis regnauit

dah. Rehoboam was forty and one years old when he began to reign, and he reigned seventeen years in Jerusalem, the city which the LORD did choose out of all the tribes of Israel, to put his name there. And his mother’s name was Naamah an Ammonitess.

in Iuda quadraginta et unius anni erat Roboam cum regnare coepisset et decem et septem annis regnauit in Hierusalem ciuitatem quam elegit Dominus ut poneret nomen suum ibi ex omnibus tribubus Israhel nomen autem matris eius Naama Ammanites 14:22 et fecit Iudas malum coram Domino 14:22 And Judah did evil in the sight of the LORD, and

they provoked him to jealousy with their sins which they had committed, above all that their fathers had done.

et inritauerunt eum super omnibus quae fecerant patres eorum in peccatis suis quae peccauerant 14:23 aedificauerunt 14:23 For they also built them high places, and images,

and groves, on every high hill, and under every green tree.

enim et ipsi sibi aras et statuas et lucos super omnem collem

excelsum et subter omnem arborem frondosam 14:24 sed 14:24 And there were also sodomites in the land: and they did according to all the abominations of the nations which the LORD cast out before the children of Israel.

et effeminati fuerunt in terra feceruntque omnes abominationes gentium quas adtriuit Dominus ante faciem filiorum Israhel 14:25 in quinto autem anno regni Roboam ascen- 14:25 And it came to pass in the fifth year of king Re-

hoboam, that Shishak king of Egypt came up against Jerusalem:

dit Sesac rex Aegypti in Hierusalem 14:26 et tulit thesau-

14:26 And he took away the treasures of the house of the LORD, and the treasures of the king’s house; he even took away all: and he took away all the shields of gold which Solomon had made.

ros domus Domini et thesauros regios et uniuersa diripuit scuta quoque aurea quae fecerat Salomon 14:27 pro quibus 14:27 And king Rehoboam made in their stead brasen

shields, and committed them unto the hands of the chief of the guard, which kept the door of the king’s house.

fecit rex Roboam scuta aerea et tradidit ea in manu ducum scutariorum et eorum qui excubabant ante ostium domus regis 14:28 cumque ingrederetur rex in domum Domini por- 14:28 And it was so, when the king went into the house

of the LORD, that the guard bare them, and brought them back into the guard chamber.

tabant ea qui praeeundi habebant officium et postea reportabant ad armamentarium scutariorum 14:29 reliqua autem 14:29 Now the rest of the acts of Rehoboam, and all that

he did, are they not written in the book of the chronicles of the kings of Judah?

sermonum Roboam et omnium quae fecit ecce scripta sunt in libro uerborum dierum regum Iuda 14:30 fuitque bellum 14:30 And there was war between Rehoboam and Jer-

oboam all their days.

inter Roboam et Hieroboam cunctis diebus 14:31 dormiuit 14:31 And Rehoboam slept with his fathers, and was

buried with his fathers in the city of David. And his mother’s name was Naamah an Ammonitess. And Abi-

itaque Roboam cum patribus suis et sepultus est cum eis in jam his son reigned in his stead. ciuitate Dauid nomen autem matris eius Naama Ammanites et regnauit Abiam filius eius pro eo

15:1 igitur in octauodecimo anno regni Hieroboam filii 15:1 Now in the eighteenth year of king Jeroboam the son

of Nebat reigned Abijam over Judah.

Nabath regnauit Abiam super Iudam 15:2 tribus annis re- 15:2 Three years reigned he in Jerusalem. and his mother’s

name was Maachah, the daughter of Abishalom.

gnauit in Hierusalem nomen matris eius Maacha filia Absalom 15:3 ambulauitque in omnibus peccatis patris sui quae 15:3 And he walked in all the sins of his father, which he

had done before him: and his heart was not perfect with the LORD his God, as the heart of David his father.

fecerat ante eum nec erat cor eius perfectum cum Domino Deo suo sicut cor Dauid patris eius 15:4 sed propter Dauid 15:4 Nevertheless for David’s sake did the LORD his God

give him a lamp in Jerusalem, to set up his son after him, and to establish Jerusalem:

dedit ei Dominus Deus suus lucernam in Hierusalem ut suscitaret filium eius post eum et staret Hierusalem 15:5 eo 15:5 Because David did that which was right in the eyes

of the LORD, and turned not aside from any thing that he commanded him all the days of his life, save only in the

quod fecisset Dauid rectum in oculis Domini et non decli- matter of Uriah the Hittite. nasset ab omnibus quae praeceperat ei cunctis diebus uitae suae excepto sermone Vriae Hetthei 15:6 attamen bellum 15:6 And there was war between Rehoboam and Jeroboam

all the days of his life.

Biblia Sacra Vulgata

fuit inter Roboam et inter Hieroboam omni tempore uitae

15:7 Now the rest of the acts of Abijam, and all that he eius 15:7 reliqua autem sermonum Abiam et omnia quae did, are they not written in the book of the chronicles of the kings of Judah? And there was war between Abijam

and Jeroboam. fecit nonne haec scripta sunt in libro uerborum dierum re-

gum Iuda fuitque proelium inter Abiam et inter Hieroboam

15:8 And Abijam slept with his fathers; and they buried 15:8 et dormiuit Abiam cum patribus suis et sepelierunt eum

him in the city of David: and Asa his son reigned in his stead.

in ciuitate Dauid regnauitque Asa filius eius pro eo 15:9

15:9 And in the twentieth year of Jeroboam king of Israel reigned Asa over Judah.

in anno ergo uicesimo Hieroboam regis Israhel regnauit Asa

15:10 And forty and one years reigned he in Jerusalem. rex Iuda 15:10 et quadraginta uno anno regnauit in Hierusa-

And his mother’s name was Maachah, the daughter of Abishalom.

lem nomen matris eius Maacha filia Absalom 15:11 et fe-

15:11 And Asa did that which was right in the eyes of the LORD, as did David his father.

cit Asa rectum ante conspectum Domini sicut Dauid pater

15:12 And he took away the sodomites out of the land, and eius 15:12 et abstulit effeminatos de terra purgauitque uni-

removed all the idols that his fathers had made.

15:13 And also Maachah his mother, even her he removed uersas sordes idolorum quae fecerant patres eius 15:13 in-

from being queen, because she had made an idol in a grove; and Asa destroyed her idol, and burnt it by the

brook Kidron. super et Maacham matrem suam amouit ne esset princeps in sacris Priapi et in luco eius quem consecrauerat subuertitque specum eius et confregit simulacrum turpissimum et con-

15:14 But the high places were not removed: nevertheless busit in torrente Cedron 15:14 excelsa autem non abstulit

Asa’s heart was perfect with the LORD all his days.

uerumtamen cor Asa perfectum erat cum Deo cunctis diebus

15:15 And he brought in the things which his father had suis 15:15 et intulit ea quae sanctificauerat pater suus et uo-

dedicated, and the things which himself had dedicated, into the house of the LORD, silver, and gold, and vessels.

uerat in domum Domini argentum et aurum et uasa 15:16

15:16 And there was war between Asa and Baasha king of Israel all their days.

bellum autem erat inter Asa et Baasa regem Israhel cunc-

15:17 And Baasha king of Israel went up against Judah, tis diebus eorum 15:17 ascendit quoque Baasa rex Israhel

and built Ramah, that he might not suffer any to go out or come in to Asa king of Judah.

in Iudam et aedificauit Rama ut non possit quispiam egredi

15:18 Then Asa took all the silver and the gold that were uel ingredi de parte Asa regis Iudae 15:18 tollens itaque

left in the treasures of the house of the LORD, and the treasures of the king’s house, and delivered them into the hand of his servants: and king Asa sent them to Benhadad, the son of Tabrimon, the son of Hezion, king of Syria, that dwelt at Damascus, saying,

Asa omne argentum et aurum quod remanserat in thesauris domus Domini et in thesauris domus regiae dedit illud in manu seruorum suorum et misit ad Benadad filium Tabremmon filii Ezion regem Syriae qui habitabat in Damasco

15:19 There is a league between me and thee, and between dicens 15:19 foedus est inter me et te et inter patrem meum

my father and thy father: behold, I have sent unto thee a present of silver and gold; come and break thy league with

Baasha king of Israel, that he may depart from me. et patrem tuum ideo misi tibi munera argentum et aurum et

peto ut uenias et irritum facias foedus quod habes cum Ba-

15:20 So Benhadad hearkened unto king Asa, and sent the asa rege Israhel et recedat a me 15:20 adquiescens Benadad

captains of the hosts which he had against the cities of Israel, and smote Ijon, and Dan, and Abelbethmaachah,

and all Cinneroth, with all the land of Naphtali. regi Asa misit principes exercitus sui in ciuitates Israhel et percusserunt Ahion et Dan et Abel domum Maacha et uniuersam Cenneroth omnem scilicet terram Nepthalim 15:21 15:21 And it came to pass, when Baasha heard thereof,

that he left off building of Ramah, and dwelt in Tirzah.

quod cum audisset Baasa intermisit aedificare Rama et reuersus est in Thersa 15:22 rex autem Asa nuntium misit in 15:22 Then king Asa made a proclamation throughout all

Judah; none was exempted: and they took away the stones of Ramah, and the timber thereof, wherewith Baasha had builded; and king Asa built with them Geba of Benjamin, and Mizpah.

omnem Iudam nemo sit excusatus et tulerunt lapides Rama et ligna eius quibus aedificauerat Baasa et extruxit de eis rex

Asa Gaba Beniamin et Maspha 15:23 reliqua autem om- 15:23 The rest of all the acts of Asa, and all his might, and all that he did, and the cities which he built, are they not written in the book of the chronicles of the kings of Judah? Nevertheless in the time of his old age he was diseased in his feet.

nium sermonum Asa et uniuersae fortitudines eius et cuncta quae fecit et ciuitates quas extruxit nonne haec scripta sunt in libro uerborum dierum regum Iuda uerumtamen in tempore senectutis suae doluit pedes 15:24 et dormiuit cum pa- 15:24 And Asa slept with his fathers, and was buried with

his fathers in the city of David his father: and Jehoshaphat his son reigned in his stead.

tribus suis et sepultus est cum eis in ciuitate Dauid patris sui regnauitque Iosaphat filius eius pro eo 15:25 Nadab uero 15:25 And Nadab the son of Jeroboam began to reign over

Israel in the second year of Asa king of Judah, and reigned over Israel two years.

filius Hieroboam regnauit super Israhel anno secundo Asa regis Iuda regnauitque super Israhel duobus annis 15:26 et 15:26 And he did evil in the sight of the LORD, and

walked in the way of his father, and in his sin wherewith he made Israel to sin.

fecit quod malum est in conspectu Domini et ambulauit in uiis patris sui et in peccatis eius quibus peccare fecit Israhel 15:27 insidiatus est autem ei Baasa filius Ahia de domo 15:27 And Baasha the son of Ahijah, of the house of Is-

sachar, conspired against him; and Baasha smote him at Gibbethon, which belonged to the Philistines; for Nadab

Isachar et percussit eum in Gebbethon quae est urbs Philis- and all Israel laid siege to Gibbethon. thinorum siquidem Nadab et omnis Israhel obsidebant Gebbethon 15:28 interfecit igitur illum Baasa in anno tertio Asa 15:28 Even in the third year of Asa king of Judah did

Baasha slay him, and reigned in his stead.

regis Iuda et regnauit pro eo 15:29 cumque regnasset per- 15:29 And it came to pass, when he reigned, that he smote

all the house of Jeroboam; he left not to Jeroboam any that breathed, until he had destroyed him, according unto the saying of the LORD, which he spake by his servant Ahijah the Shilonite:

cussit omnem domum Hieroboam non dimisit ne unam quidem animam de semine eius donec deleret eum iuxta uerbum Domini quod locutus fuerat in manu serui sui Ahiae Silonitis 15:30 propter peccata Hieroboam quae peccauerat et 15:30 Because of the sins of Jeroboam which he sinned,

and which he made Israel sin, by his provocation wherewith he provoked the LORD God of Israel to anger.

quibus peccare fecerat Israhel et propter delictum quo inritauerat Dominum Deum Israhel 15:31 reliqua autem sermo- 15:31 Now the rest of the acts of Nadab, and all that he

did, are they not written in the book of the chronicles of the kings of Israel?

num Nadab et omnia quae operatus est nonne haec scripta sunt in libro uerborum dierum regum Israhel 15:32 fuitque 15:32 And there was war between Asa and Baasha king of

Israel all their days.

bellum inter Asa et Baasa regem Israhel cunctis diebus eorum 15:33 anno tertio Asa regis Iuda regnauit Baasa filius 15:33 In the third year of Asa king of Judah began Baasha

the son of Ahijah to reign over all Israel in Tirzah, twenty and four years.

Ahia super omnem Israhel in Thersa uiginti quattuor annis 15:34 et fecit malum coram Domino ambulauitque in uia Hie- 15:34 And he did evil in the sight of the LORD, and

walked in the way of Jeroboam, and in his sin wherewith he made Israel to sin.

Biblia Sacra Vulgata

roboam et in peccatis eius quibus peccare fecit Israhel

16:1 Then the word of the LORD came to Jehu the son of 16:1 factus est autem sermo Domini ad Hieu filium Anani

Hanani against Baasha, saying,

16:2 Forasmuch as I exalted thee out of the dust, and made contra Baasa dicens 16:2 pro eo quod exaltaui te de puluere

thee prince over my people Israel; and thou hast walked in the way of Jeroboam, and hast made my people Israel to

sin, to provoke me to anger with their sins; et posui ducem super populum meum Israhel tu autem am-

bulasti in uia Hieroboam et peccare fecisti populum meum

16:3 Behold, I will take away the posterity of Baasha, and Israhel ut me inritares in peccatis eorum 16:3 ecce ego de-

the posterity of his house; and will make thy house like the house of Jeroboam the son of Nebat.

metam posteriora Baasa et posteriora domus eius et faciam

16:4 Him that dieth of Baasha in the city shall the dogs domum tuam sicut domum Hieroboam filii Nabath 16:4 qui eat; and him that dieth of his in the fields shall the fowls of the air eat.

mortuus fuerit de Baasa in ciuitate comedent eum canes et qui mortuus fuerit ex eo in regione comedent eum uolucres

16:5 Now the rest of the acts of Baasha, and what he did, caeli 16:5 reliqua autem sermonum Baasa et quaecumque

and his might, are they not written in the book of the chronicles of the kings of Israel?

fecit et proelia eius nonne haec scripta sunt in libro uerbo-

16:6 So Baasha slept with his fathers, and was buried in rum dierum regum Israhel 16:6 dormiuit ergo Baasa cum Tirzah: and Elah his son reigned in his stead.

patribus suis sepultusque est in Thersa et regnauit Hela fi-

16:7 And also by the hand of the prophet Jehu the son of lius eius pro eo 16:7 cum autem in manu Hieu filii Anani Hanani came the word of the LORD against Baasha, and against his house, even for all the evil that he did in the sight of the LORD, in provoking him to anger with the work of his hands, in being like the house of Jeroboam; and because he killed him.

prophetae uerbum Domini factum esset contra Baasa et contra domum eius et contra omne malum quod fecerat coram Domino ad inritandum eum in operibus manuum suarum ut fieret sicut domus Hieroboam ob hanc causam occidit eum

16:8 In the twenty and sixth year of Asa king of Judah 16:8 anno uicesimo sexto Asa regis Iuda regnauit Hela filius

began Elah the son of Baasha to reign over Israel in Tirzah, two years.

Baasa super Israhel in Thersa duobus annis 16:9 et rebella-

16:9 And his servant Zimri, captain of half his chariots, conspired against him, as he was in Tirzah, drinking himself drunk in the house of Arza steward of his house in Tirzah.

uit contra eum seruus suus Zamri dux mediae partis equitum erat autem Hela in Thersa bibens et temulentus in domo Arsa

16:10 And Zimri went in and smote him, and killed him, praefecti Thersa 16:10 inruens ergo Zamri percussit et oc-

in the twenty and seventh year of Asa king of Judah, and reigned in his stead.

cidit eum anno uicesimo septimo Asa regis Iuda et regnauit

16:11 And it came to pass, when he began to reign, as pro eo 16:11 cumque regnasset et sedisset super solium eius

soon as he sat on his throne, that he slew all the house of Baasha: he left him not one that pisseth against a wall,

neither of his kinsfolks, nor of his friends. percussit omnem domum Baasa et non dereliquit ex eo min-16:12 Thus did Zimri destroy all the house of Baasha, ac- gentem ad parietem et propinquos et amicos eius 16:12 de-

cording to the word of the LORD, which he spake against Baasha by Jehu the prophet.

leuitque Zamri omnem domum Baasa iuxta uerbum Domini

16:13 For all the sins of Baasha, and the sins of Elah his quod locutus fuerat ad Baasa in manu Hieu prophetae 16:13

son, by which they sinned, and by which they made Israel to sin, in provoking the LORD God of Israel to anger with

their vanities. propter uniuersa peccata Baasa et peccata Hela filii eius qui

peccauerunt et peccare fecerunt Israhel prouocantes Domi-

16:14 Now the rest of the acts of Elah, and all that he did, num Deum Israhel in uanitatibus suis 16:14 reliqua autem

are they not written in the book of the chronicles of the kings of Israel?

sermonum Hela et omnia quae fecit nonne haec scripta sunt in libro uerborum dierum regum Israhel 16:15 anno uice- 16:15 In the twenty and seventh year of Asa king of Ju-

dah did Zimri reign seven days in Tirzah. And the people were encamped against Gibbethon, which belonged to the

simo et septimo Asa regis Iuda regnauit Zamri septem die- Philistines. bus in Thersa porro exercitus obsidebat Gebbethon urbem Philisthinorum 16:16 cumque audisset rebellasse Zamri et 16:16 And the people that were encamped heard say, Zimri

hath conspired, and hath also slain the king: wherefore all Israel made Omri, the captain of the host, king over Israel

occidisse regem fecit sibi regem omnis Israhel Amri qui erat that day in the camp. princeps militiae super Israhel in die illa in castris 16:17 16:17 And Omri went up from Gibbethon, and all Israel

with him, and they besieged Tirzah.

ascendit ergo Amri et omnis Israhel cum eo de Gebbethon et obsidebant Thersa 16:18 uidens autem Zamri quod ex- 16:18 And it came to pass, when Zimri saw that the city

was taken, that he went into the palace of the king’s house, and burnt the king’s house over him with fire, and died.

pugnanda esset ciuitas ingressus est palatium et succendit secum domum regiam et mortuus est 16:19 in peccatis suis 16:19 For his sins which he sinned in doing evil in the

sight of the LORD, in walking in the way of Jeroboam, and in his sin which he did, to make Israel to sin.

quae peccauerat faciens malum coram Domino et ambulans in uia Hieroboam et in peccato eius quo fecit peccare Israhel 16:20 reliqua autem sermonum Zamri et insidiarum 16:20 Now the rest of the acts of Zimri, and his treason

that he wrought, are they not written in the book of the chronicles of the kings of Israel?

eius et tyrannidis nonne haec scripta sunt in libro uerborum dierum regum Israhel 16:21 tunc diuisus est populus Isra- 16:21 Then were the people of Israel divided into two

parts: half of the people followed Tibni the son of Ginath, to make him king; and half followed Omri.

hel in duas partes media pars populi sequebatur Thebni filium Gineth ut constitueret eum regem et media pars Amri

16:22 praeualuit autem populus qui erat cum Amri populo 16:22 But the people that followed Omri prevailed against the people that followed Tibni the son of Ginath: so Tibni died, and Omri reigned.

qui sequebatur Thebni filium Gineth mortuusque est Thebni et regnauit Amri 16:23 anno tricesimo primo Asa regis Iuda 16:23 In the thirty and first year of Asa king of Judah

began Omri to reign over Israel, twelve years: six years reigned he in Tirzah.

regnauit Amri super Israhel duodecim annis in Thersa regnauit sex annis 16:24 emitque montem Samariae a Somer 16:24 And he bought the hill Samaria of Shemer for two

talents of silver, and built on the hill, and called the name of the city which he built, after the name of Shemer, owner

duobus talentis argenti et aedificauit eam et uocauit nomen of the hill, Samaria. ciuitatis quam extruxerat nomine Somer domini montis Samariae 16:25 fecit autem Amri malum in conspectu Domini 16:25 But Omri wrought evil in the eyes of the LORD, and

did worse than all that were before him.

et operatus est nequiter super omnes qui fuerant ante eum 16:26 ambulauitque in omni uia Hieroboam filii Nabath et in 16:26 For he walked in all the way of Jeroboam the son

of Nebat, and in his sin wherewith he made Israel to sin, to provoke the LORD God of Israel to anger with their

peccatis eius quibus peccare fecerat Israhel ut inritaret Do- vanities. minum Deum Israhel in uanitatibus suis 16:27 reliqua au- 16:27 Now the rest of the acts of Omri which he did, and

his might that he shewed, are they not written in the book of the chronicles of the kings of Israel?

tem sermonum Amri et proelia eius quae gessit nonne haec scripta sunt in libro uerborum dierum regum Israhel 16:28 16:28 So Omri slept with his fathers, and was buried in

Samaria: and Ahab his son reigned in his stead.

et dormiuit Amri cum patribus suis et sepultus est in Samaria

Biblia Sacra Vulgata

16:29 And in the thirty and eighth year of Asa king of regnauitque Ahab filius eius pro eo 16:29 Ahab uero filius

Judah began Ahab the son of Omri to reign over Israel: and Ahab the son of Omri reigned over Israel in Samaria

twenty and two years. Amri regnauit super Israhel anno tricesimo octauo Asa regis

Iuda et regnauit Ahab filius Amri super Israhel in Samaria

16:30 And Ahab the son of Omri did evil in the sight of uiginti et duobus annis 16:30 et fecit Ahab filius Amri ma-

the LORD above all that were before him.

lum in conspectu Domini super omnes qui fuerunt ante eum

16:31 And it came to pass, as if it had been a light thing 16:31 nec suffecit ei ut ambularet in peccatis Hieroboam filii

for him to walk in the sins of Jeroboam the son of Nebat, that he took to wife Jezebel the daughter of Ethbaal king of the Zidonians, and went and served Baal, and worshipped him.

Nabath insuper duxit uxorem Hiezabel filiam Ethbaal regis

16:32 And he reared up an altar for Baal in the house of Sidoniorum et abiit et seruiuit Baal et adorauit eum 16:32

Baal, which he had built in Samaria.

et posuit aram Baal in templo Baal quod aedificauerat in Sa-

16:33 And Ahab made a grove; and Ahab did more to pro- maria 16:33 et plantauit lucum et addidit Ahab in opere suo

voke the LORD God of Israel to anger than all the kings of Israel that were before him.

inritans Dominum Deum Israhel super omnes reges Israhel

16:34 In his days did Hiel the Bethelite build Jericho: he qui fuerant ante eum 16:34 in diebus eius aedificauit Ahiel

laid the foundation thereof in Abiram his firstborn, and set up the gates thereof in his youngest son Segub, according to the word of the LORD, which he spake by Joshua the son of Nun.

de Bethel Hiericho in Abiram primitiuo suo fundauit eam et in Segub nouissimo suo posuit portas eius iuxta uerbum Domini quod locutus fuerat in manu Iosue filii Nun

17:1 And Elijah the Tishbite, who was of the inhabitants 17:1 et dixit Helias Thesbites de habitatoribus Galaad ad

of Gilead, said unto Ahab, As the LORD God of Israel liveth, before whom I stand, there shall not be dew nor

rain these years, but according to my word. Ahab uiuit Dominus Deus Israhel in cuius conspectu sto si 17:2 And the word of the LORD came unto him, saying, erit annis his ros et pluuia nisi iuxta oris mei uerba 17:2

17:3 Get thee hence, and turn thee eastward, and hide thy- et factum est uerbum Domini ad eum dicens 17:3 recede self by the brook Cherith, that is before Jordan.

hinc et uade contra orientem et abscondere in torrente Cha-

17:4 And it shall be, that thou shalt drink of the brook; and rith qui est contra Iordanem 17:4 et ibi de torrente bibes I have commanded the ravens to feed thee there.

17:5 So he went and did according unto the word of the coruisque praecepi ut pascant te ibi 17:5 abiit ergo et fecit

LORD: for he went and dwelt by the brook Cherith, that is before Jordan.

iuxta uerbum Domini cumque abisset sedit in torrente Cha-

17:6 And the ravens brought him bread and flesh in the rith qui est contra Iordanem 17:6 corui quoque deferebant

morning, and bread and flesh in the evening; and he drank of the brook.

panem et carnes mane similiter panem et carnes uesperi et

17:7 And it came to pass after a while, that the brook dried bibebat de torrente 17:7 post dies autem siccatus est torrens

up, because there had been no rain in the land.

17:8 And the word of the LORD came unto him, saying, non enim pluerat super terram 17:8 factus est igitur sermo

17:9 Arise, get thee to Zarephath, which belongeth to Zi- Domini ad eum dicens 17:9 surge et uade in Sareptha Sidon, and dwell there: behold, I have commanded a widow woman there to sustain thee.

doniorum et manebis ibi praecepi enim ibi mulieri uiduae ut

17:10 So he arose and went to Zarephath. And when he pascat te 17:10 surrexit et abiit Sareptham cumque uenisset

came to the gate of the city, behold, the widow woman was there gathering of sticks: and he called to her, and said, Fetch me, I pray thee, a little water in a vessel, that I may drink.

ad portam ciuitatis apparuit ei mulier uidua colligens ligna et uocauit eam dixitque da mihi paululum aquae in uase ut

17:11 And as she was going to fetch it, he called to her, bibam 17:11 cumque illa pergeret ut adferret clamauit post

and said, Bring me, I pray thee, a morsel of bread in thine hand.

tergum eius dicens adfer mihi obsecro et buccellam panis in manu tua 17:12 quae respondit uiuit Dominus Deus tuus 17:12 And she said, As the LORD thy God liveth, I have

not a cake, but an handful of meal in a barrel, and a little oil in a cruse: and, behold, I am gathering two sticks, that I may go in and dress it for me and my son, that we may eat it, and die.

quia non habeo panem nisi quantum pugillus capere potest farinae in hydria et paululum olei in lecytho en colligo duo ligna ut ingrediar et faciam illud mihi et filio meo ut come-

damus et moriamur 17:13 ad quam Helias ait noli timere 17:13 And Elijah said unto her, Fear not; go and do as thou hast said: but make me thereof a little cake first, and bring it unto me, and after make for thee and for thy son.

sed uade et fac sicut dixisti uerumtamen mihi primum fac de ipsa farinula subcinericium panem paruulum et adfer ad me tibi autem et filio tuo facies postea 17:14 haec autem dicit 17:14 For thus saith the LORD God of Israel, The barrel

of meal shall not waste, neither shall the cruse of oil fail, until the day that the LORD sendeth rain upon the earth.

Dominus Deus Israhel hydria farinae non deficiet nec lecythus olei minuetur usque ad diem in qua daturus est Dominus pluuiam super faciem terrae 17:15 quae abiit et fecit iuxta 17:15 And she went and did according to the saying of

Elijah: and she, and he, and her house, did eat many days.

uerbum Heliae et comedit ipse et illa et domus eius et ex illa die 17:16 hydria farinae non defecit et lecythus olei non est 17:16 And the barrel of meal wasted not, neither did the

cruse of oil fail, according to the word of the LORD, which he spake by Elijah.

inminutus iuxta uerbum Domini quod locutus fuerat in manu Heliae 17:17 factum est autem post uerba haec aegrotauit 17:17 And it came to pass after these things, that the son

of the woman, the mistress of the house, fell sick; and his sickness was so sore, that there was no breath left in him.

filius mulieris matris familiae et erat languor fortis nimis ita ut non remaneret in eo halitus 17:18 dixit ergo ad Heliam 17:18 And she said unto Elijah, What have I to do with

thee, O thou man of God? art thou come unto me to call my sin to remembrance, and to slay my son?

quid mihi et tibi uir Dei ingressus es ad me ut rememorarentur iniquitates meae et interficeres filium meum 17:19 et 17:19 And he said unto her, Give me thy son. And he took

him out of her bosom, and carried him up into a loft, where he abode, and laid him upon his own bed.

ait ad eam da mihi filium tuum tulitque eum de sinu illius et portauit in cenaculum ubi ipse manebat et posuit super lectulum suum 17:20 et clamauit ad Dominum et dixit Domine 17:20 And he cried unto the LORD, and said, O LORD

my God, hast thou also brought evil upon the widow with whom I sojourn, by slaying her son?

Deus meus etiamne uiduam apud quam ego utcumque sustentor adflixisti ut interficeres filium eius 17:21 et expandit 17:21 And he stretched himself upon the child three times,

and cried unto the LORD, and said, O LORD my God, I pray thee, let this child’s soul come into him again.

se atque mensus est super puerum tribus uicibus clamauitque ad Dominum et ait Domine Deus meus reuertatur oro anima pueri huius in uiscera eius 17:22 exaudiuit Dominus uocem 17:22 And the LORD heard the voice of Elijah; and the

soul of the child came into him again, and he revived.

Heliae et reuersa est anima pueri intra eum et reuixit 17:23 17:23 And Elijah took the child, and brought him down

out of the chamber into the house, and delivered him unto his mother: and Elijah said, See, thy son liveth.

tulitque Helias puerum et deposuit eum de cenaculo in inferiorem domum et tradidit matri suae et ait illi en uiuit filius tuus 17:24 dixitque mulier ad Heliam nunc in isto cognoui 17:24 And the woman said to Elijah, Now by this I know

that thou art a man of God, and that the word of the LORD in thy mouth is truth.

quoniam uir Dei es tu et uerbum Domini in ore tuo uerum

Biblia Sacra Vulgata

est

18:1 And it came to pass after many days, that the word of 18:1 post dies multos uerbum Domini factum est ad Hethe LORD came to Elijah in the third year, saying, Go, shew thyself unto Ahab; and I will send rain upon the

earth. liam in anno tertio dicens uade et ostende te Ahab ut dem 18:2 And Elijah went to shew himself unto Ahab. And pluuiam super faciem terrae 18:2 iuit ergo Helias ut os-

there was a sore famine in Samaria.

tenderet se Ahab erat autem fames uehemens in Samaria

18:3 And Ahab called Obadiah, which was the governor 18:3 uocauitque Ahab Abdiam dispensatorem domus suae

of his house. (Now Obadiah feared the LORD greatly: 18:4 For it was so, when Jezebel cut off the prophets of Abdias autem timebat Dominum ualde 18:4 nam cum inthe LORD, that Obadiah took an hundred prophets, and hid them by fifty in a cave, and fed them with bread and

water.) terficeret Hiezabel prophetas Domini tulit ille centum pro-

phetas et abscondit eos quinquagenos in speluncis et pauit

18:5 And Ahab said unto Obadiah, Go into the land, unto eos pane et aqua 18:5 dixit ergo Ahab ad Abdiam uade in all fountains of water, and unto all brooks: peradventure we may find grass to save the horses and mules alive, that

we lose not all the beasts. terram ad uniuersos fontes aquarum et in cunctas ualles si

forte inuenire possimus herbam et saluare equos et mulos

18:6 So they divided the land between them to pass et non penitus iumenta intereant 18:6 diuiseruntque sibi re-

throughout it: Ahab went one way by himself, and Obadiah went another way by himself.

giones ut circuirent eas Ahab ibat per uiam unam et Abdias

18:7 And as Obadiah was in the way, behold, Elijah met per uiam alteram seorsum 18:7 cumque esset Abdias in uia

him: and he knew him, and fell on his face, and said, Art thou that my lord Elijah?

Helias occurrit ei qui cum cognouisset eum cecidit super

18:8 And he answered him, I am: go, tell thy lord, Behold, faciem suam et ait num tu es domine mi Helias 18:8 cui Elijah is here.

18:9 And he said, What have I sinned, that thou wouldest ille respondit ego uade dic domino tuo adest Helias 18:9 et

deliver thy servant into the hand of Ahab, to slay me?

ille quid peccaui inquit quoniam trades me seruum tuum in

18:10 As the LORD thy God liveth, there is no nation or manu Ahab ut interficiat me 18:10 uiuit Dominus Deus tuus

kingdom, whither my lord hath not sent to seek thee: and when they said, He is not there; he took an oath of the

kingdom and nation, that they found thee not. non est gens aut regnum quo non miserit dominus meus te

requirens et respondentibus cunctis non est hic adiurauit re-

18:11 And now thou sayest, Go, tell thy lord, Behold, Eli- gna singula et gentes eo quod minime repperireris 18:11 et

jah is here.

18:12 And it shall come to pass, as soon as I am gone from nunc dicis mihi uade et dic domino tuo adest Helias 18:12

thee, that the Spirit of the LORD shall carry thee whither I know not; and so when I come and tell Ahab, and he cannot find thee, he shall slay me: but I thy servant fear the LORD from my youth.

cumque recessero a te spiritus Domini asportabit te in locum quem ego ignoro ingressus nuntiabo Ahab et non inueniet te et interficiet me seruus autem tuus timet Dominum ab

18:13 Was it not told my lord what I did when Jezebel slew infantia sua 18:13 numquid non indicatum est tibi domino

the prophets of the LORD, how I hid an hundred men of the LORD’s prophets by fifty in a cave, and fed them with

bread and water? meo quid fecerim cum interficeret Hiezabel prophetas Domini quod absconderim de prophetis Domini centum uiros quinquagenos et quinquagenos in speluncis et pauerim eos

18:14 And now thou sayest, Go, tell thy lord, Behold, Eli- pane et aqua 18:14 et nunc tu dicis uade et dic domino tuo jah is here: and he shall slay me.

18:15 And Elijah said, As the LORD of hosts liveth, be- adest Helias ut interficiat me 18:15 dixit Helias uiuit Domi-

fore whom I stand, I will surely shew myself unto him to day.

nus exercituum ante cuius uultum sto quia hodie apparebo ei 18:16 abiit ergo Abdias in occursum Ahab et indicauit 18:16 So Obadiah went to meet Ahab, and told him: and

Ahab went to meet Elijah.

ei uenitque Ahab in occursum Heliae 18:17 et cum uidis- 18:17 And it came to pass, when Ahab saw Elijah, that

Ahab said unto him, Art thou he that troubleth Israel?

set eum ait tune es ille qui conturbas Israhel 18:18 et ille ait 18:18 And he answered, I have not troubled Israel; but

thou, and thy father’s house, in that ye have forsaken the commandments of the LORD, and thou hast followed

non turbaui Israhel sed tu et domus patris tui qui dereliquistis Baalim.

mandata Domini et secuti estis Baalim 18:19 uerumtamen 18:19 Now therefore send, and gather to me all Israel unto mount Carmel, and the prophets of Baal four hundred and fifty, and the prophets of the groves four hundred, which

nunc mitte et congrega ad me uniuersum Israhel in monte eat at Jezebel’s table. Carmeli et prophetas Baal quadringentos quinquaginta prophetasque lucorum quadringentos qui comedunt de mensa Hiezabel 18:20 misit Ahab ad omnes filios Israhel et con- 18:20 So Ahab sent unto all the children of Israel, and

gathered the prophets together unto mount Carmel.

gregauit prophetas in monte Carmeli 18:21 accedens autem 18:21 And Elijah came unto all the people, and said, How

long halt ye between two opinions? if the LORD be God, follow him: but if Baal, then follow him. And the people

Helias ad omnem populum ait usquequo claudicatis in duas answered him not a word. partes si Dominus est Deus sequimini eum si autem Baal sequimini illum et non respondit ei populus uerbum 18:22 18:22 Then said Elijah unto the people, I, even I only, re-

main a prophet of the LORD; but Baal’s prophets are four hundred and fifty men.

et ait rursum Helias ad populum ego remansi propheta Domini solus prophetae autem Baal quadringenti et quinquaginta uiri sunt 18:23 dentur nobis duo boues et illi eligant 18:23 Let them therefore give us two bullocks; and let

them choose one bullock for themselves, and cut it in pieces, and lay it on wood, and put no fire under: and I will dress the other bullock, and lay it on wood, and put no fire under:

bouem unum et in frusta caedentes ponant super ligna ignem autem non subponant et ego faciam bouem alterum et inponam super ligna ignemque non subponam 18:24 inuo- 18:24 And call ye on the name of your gods, and I will call

on the name of the LORD: and the God that answereth by fire, let him be God. And all the people answered and said,

cate nomina deorum uestrorum et ego inuocabo nomen Do- It is well spoken. mini et deus qui exaudierit per ignem ipse sit Deus respondens omnis populus ait optima propositio 18:25 dixit ergo 18:25 And Elijah said unto the prophets of Baal, Choose

you one bullock for yourselves, and dress it first; for ye are many; and call on the name of your gods, but put no

Helias prophetis Baal eligite uobis bouem unum et facite fire under. primi quia uos plures estis et inuocate nomina deorum uestrorum ignemque non subponatis 18:26 qui cum tulissent 18:26 And they took the bullock which was given them,

and they dressed it, and called on the name of Baal from morning even until noon, saying, O Baal, hear us. But there was no voice, nor any that answered. And they leaped upon the altar which was made.

bouem quem dederat eis fecerunt et inuocabant nomen Baal de mane usque ad meridiem dicentes Baal exaudi nos et non erat uox nec qui responderet transiliebantque altare quod fecerant 18:27 cumque esset iam meridies inludebat eis He- 18:27 And it came to pass at noon, that Elijah mocked

them, and said, Cry aloud: for he is a god; either he is talking, or he is pursuing, or he is in a journey, or perad-

lias dicens clamate uoce maiore deus enim est et forsitan venture he sleepeth, and must be awaked. loquitur aut in diuersorio est aut in itinere aut certe dormit ut excitetur 18:28 clamabant ergo uoce magna et incidebant 18:28 And they cried aloud, and cut themselves after their

manner with knives and lancets, till the blood gushed out upon them.

Biblia Sacra Vulgata

se iuxta ritum suum cultris et lanceolis donec perfunderen-

18:29 And it came to pass, when midday was past, and tur sanguine 18:29 postquam autem transiit meridies et illis

they prophesied until the time of the offering of the evening sacrifice, that there was neither voice, nor any to

answer, nor any that regarded. prophetantibus uenerat tempus quo sacrificium offerri solet

nec audiebatur uox neque aliquis respondebat nec adtende-

18:30 And Elijah said unto all the people, Come near unto bat orantes 18:30 dixit Helias omni populo uenite ad me et

me. And all the people came near unto him. And he repaired the altar of the LORD that was broken down.

accedente ad se populo curauit altare Domini quod destruc-

18:31 And Elijah took twelve stones, according to the tum fuerat 18:31 et tulit duodecim lapides iuxta numerum

number of the tribes of the sons of Jacob, unto whom the word of the LORD came, saying, Israel shall be thy name:

tribuum filiorum Iacob ad quem factus est sermo Domini

18:32 And with the stones he built an altar in the name of dicens Israhel erit nomen tuum 18:32 et aedificauit lapidi-

the LORD: and he made a trench about the altar, as great as would contain two measures of seed.

bus altare in nomine Domini fecitque aquaeductum quasi per

18:33 And he put the wood in order, and cut the bullock duas aratiunculas in circuitu altaris 18:33 et conposuit ligna

in pieces, and laid him on the wood, and said, Fill four barrels with water, and pour it on the burnt sacrifice, and

on the wood. diuisitque per membra bouem et posuit super ligna 18:34 et

18:34 And he said, Do it the second time. And they did it the second time. And he said, Do it the third time. And they did it the third time.

ait implete quattuor hydrias aqua et fundite super holocaustum et super ligna rursumque dixit etiam secundo hoc facite qui cum fecissent et secundo ait etiam tertio id ipsum facite

18:35 And the water ran round about the altar; and he filled feceruntque et tertio 18:35 et currebant aquae circa altare

the trench also with water.

18:36 And it came to pass at the time of the offering of the et fossa aquaeductus repleta est 18:36 cumque iam tempus

evening sacrifice, that Elijah the prophet came near, and said, LORD God of Abraham, Isaac, and of Israel, let it be known this day that thou art God in Israel, and that I am thy servant, and that I have done all these things at thy word.

esset ut offerretur holocaustum accedens Helias propheta ait Domine Deus Abraham Isaac et Israhel hodie ostende quia tu es Deus Israhel et ego seruus tuus et iuxta praeceptum

18:37 Hear me, O LORD, hear me, that this people may tuum feci omnia uerba haec 18:37 exaudi me Domine ex-

know that thou art the LORD God, and that thou hast turned their heart back again.

audi me ut discat populus iste quia tu es Dominus Deus et

18:38 Then the fire of the LORD fell, and consumed the tu conuertisti cor eorum iterum 18:38 cecidit autem ignis

burnt sacrifice, and the wood, and the stones, and the dust, and licked up the water that was in the trench.

Domini et uorauit holocaustum et ligna et lapides puluerem

18:39 And when all the people saw it, they fell on their quoque et aquam quae erat in aquaeductu lambens 18:39 faces: and they said, The LORD, he is the God; the LORD, he is the God.

quod cum uidisset omnis populus cecidit in faciem suam et

18:40 And Elijah said unto them, Take the prophets of ait Dominus ipse est Deus Dominus ipse est Deus 18:40 Baal; let not one of them escape. And they took them: and Elijah brought them down to the brook Kishon, and

slew them there. dixitque Helias ad eos adprehendite prophetas Baal et ne

unus quidem fugiat ex eis quos cum conprehendissent duxit

18:41 And Elijah said unto Ahab, Get thee up, eat and eos Helias ad torrentem Cison et interfecit eos ibi 18:41 et

drink; for there is a sound of abundance of rain.

ait Helias ad Ahab ascende comede et bibe quia sonus mul-

18:42 So Ahab went up to eat and to drink. And Elijah tae pluuiae est 18:42 ascendit Ahab ut comederet et biberet

went up to the top of Carmel; and he cast himself down upon the earth, and put his face between his knees,

Helias autem ascendit in uertice Carmeli et pronus in terram posuit faciem inter genua sua 18:43 et dixit ad puerum 18:43 And said to his servant, Go up now, look toward

the sea. And he went up, and looked, and said, There is nothing. And he said, Go again seven times.

suum ascende et prospice contra mare qui cum ascendisset et contemplatus esset ait non est quicquam et rursum ait illi reuertere septem uicibus 18:44 in septima autem uice ecce 18:44 And it came to pass at the seventh time, that he said,

Behold, there ariseth a little cloud out of the sea, like a man’s hand. And he said, Go up, say unto Ahab, Prepare

nubicula parua quasi uestigium hominis ascendebat de mari thy chariot, and get thee down that the rain stop thee not. qui ait ascende et dic Ahab iunge et descende ne occupet te pluuia 18:45 cumque se uerterent huc atque illuc ecce 18:45 And it came to pass in the mean while, that the

heaven was black with clouds and wind, and there was a great rain. And Ahab rode, and went to Jezreel.

caeli contenebrati sunt et nubes et uentus et facta est pluuia grandis ascendens itaque Ahab abiit in Hiezrahel 18:46 et 18:46 And the hand of the LORD was on Elijah; and he

girded up his loins, and ran before Ahab to the entrance of Jezreel.

manus Domini facta est super Heliam accinctisque lumbis currebat ante Ahab donec ueniret in Hiezrahel

19:1 nuntiauit autem Ahab Hiezabel omnia quae fecerat 19:1 And Ahab told Jezebel all that Elijah had done, and

withal how he had slain all the prophets with the sword.

Helias et quomodo occidisset uniuersos prophetas gladio 19:2 misitque Hiezabel nuntium ad Heliam dicens haec mihi 19:2 Then Jezebel sent a messenger unto Elijah, saying,

So let the gods do to me, and more also, if I make not thy life as the life of one of them by to morrow about this time.

faciant dii et haec addant nisi hac hora cras posuero animam tuam sicut animam unius ex illis 19:3 timuit ergo Helias et 19:3 And when he saw that, he arose, and went for his life,

and came to Beersheba, which belongeth to Judah, and left his servant there.

surgens abiit quocumque eum ferebat uoluntas uenitque in Bersabee Iuda et dimisit ibi puerum suum 19:4 et perrexit 19:4 But he himself went a day’s journey into the wilder-

ness, and came and sat down under a juniper tree: and he requested for himself that he might die; and said, It is enough; now, O LORD, take away my life; for I am not better than my fathers.

in desertum uia unius diei cumque uenisset et sederet subter unam iuniperum petiuit animae suae ut moreretur et ait sufficit mihi Domine tolle animam meam neque enim melior sum quam patres mei 19:5 proiecitque se et obdormi- 19:5 And as he lay and slept under a juniper tree, behold,

then an angel touched him, and said unto him, Arise and eat.

uit in umbra iuniperi et ecce angelus tetigit eum et dixit illi surge comede 19:6 respexit et ecce ad caput suum subci- 19:6 And he looked, and, behold, there was a cake baken

on the coals, and a cruse of water at his head. And he did eat and drink, and laid him down again.

nericius panis et uas aquae comedit ergo et bibit et rursum obdormiuit 19:7 reuersusque est angelus Domini secundo 19:7 And the angel of the LORD came again the second

time, and touched him, and said, Arise and eat; because the journey is too great for thee.

et tetigit eum dixitque illi surge comede grandis enim tibi restat uia 19:8 qui cum surrexisset comedit et bibit et am- 19:8 And he arose, and did eat and drink, and went in

the strength of that meat forty days and forty nights unto Horeb the mount of God.

bulauit in fortitudine cibi illius quadraginta diebus et quadraginta noctibus usque ad montem Dei Horeb 19:9 cumque 19:9 And he came thither unto a cave, and lodged there;

and, behold, the word of the LORD came to him, and he said unto him, What doest thou here, Elijah?

uenisset illuc mansit in spelunca et ecce sermo Domini ad eum dixitque illi quid hic agis Helia 19:10 at ille respondit 19:10 And he said, I have been very jealous for the LORD

God of hosts: for the children of Israel have forsaken thy covenant, thrown down thine altars, and slain thy prophets with the sword; and I, even I only, am left; and they seek my life, to take it away.

zelo zelatus sum pro Domino Deo exercituum quia dereli-

Biblia Sacra Vulgata

querunt pactum Domini filii Israhel altaria tua destruxerunt et prophetas tuos occiderunt gladio et derelictus sum ego so-

19:11 And he said, Go forth, and stand upon the mount lus et quaerunt animam meam ut auferant eam 19:11 et ait

before the LORD. And, behold, the LORD passed by, and a great and strong wind rent the mountains, and brake in pieces the rocks before the LORD; but the LORD was not in the wind: and after the wind an earthquake; but the LORD was not in the earthquake:

ei egredere et sta in monte coram Domino et ecce Dominus transit et spiritus grandis et fortis subuertens montes et conterens petras ante Dominum non in spiritu Dominus et post

19:12 And after the earthquake a fire; but the LORD was spiritum commotio non in commotione Dominus 19:12 et

not in the fire: and after the fire a still small voice.

post commotionem ignis non in igne Dominus et post ig-

19:13 And it was so, when Elijah heard it, that he wrapped nem sibilus aurae tenuis 19:13 quod cum audisset Helias his face in his mantle, and went out, and stood in the entering in of the cave. And, behold, there came a voice unto

him, and said, What doest thou here, Elijah? operuit uultum suum pallio et egressus stetit in ostio spelun-19:14 And he said, I have been very jealous for the LORD cae et ecce uox ad eum dicens quid agis hic Helia 19:14 et

God of hosts: because the children of Israel have forsaken thy covenant, thrown down thine altars, and slain thy prophets with the sword; and I, even I only, am left; and they seek my life, to take it away.

ille respondit zelo zelatus sum pro Domino Deo exercituum quia dereliquerunt pactum tuum filii Israhel altaria tua destruxerunt et prophetas tuos occiderunt gladio et derelictus sum ego solus et quaerunt animam meam ut auferant eam

19:15 And the LORD said unto him, Go, return on thy way 19:15 et ait Dominus ad eum uade et reuertere in uiam tuam

to the wilderness of Damascus: and when thou comest, anoint Hazael to be king over Syria:

per desertum in Damascum cumque perueneris ungues Aza-

19:16 And Jehu the son of Nimshi shalt thou anoint to be hel regem super Syriam 19:16 et Hieu filium Namsi ungues

king over Israel: and Elisha the son of Shaphat of Abelmeholah shalt thou anoint to be prophet in thy room.

regem super Israhel Heliseum autem filium Saphat qui est de

19:17 And it shall come to pass, that him that escapeth the Abelmaula ungues prophetam pro te 19:17 et erit quicumsword of Hazael shall Jehu slay: and him that escapeth from the sword of Jehu shall Elisha slay.

que fugerit gladium Azahel occidet eum Hieu et qui fuge-

19:18 Yet I have left me seven thousand in Israel, all the rit gladium Hieu interficiet eum Heliseus 19:18 et derelin-

knees which have not bowed unto Baal, and every mouth which hath not kissed him.

quam mihi in Israhel septem milia uniuersorum genua quae non sunt incuruata Baal et omne os quod non adorauit eum

19:19 So he departed thence, and found Elisha the son of osculans manum 19:19 profectus ergo inde repperit Heli-

Shaphat, who was plowing with twelve yoke of oxen before him, and he with the twelfth: and Elijah passed by

him, and cast his mantle upon him. seum filium Saphat arantem duodecim iugis boum et ipse

in duodecim arantibus unus erat cumque uenisset Helias ad

19:20 And he left the oxen, and ran after Elijah, and said, eum misit pallium suum super illum 19:20 qui statim relic-

Let me, I pray thee, kiss my father and my mother, and then I will follow thee. And he said unto him, Go back

again: for what have I done to thee? tis bubus cucurrit post Heliam et ait osculer oro te patrem

meum et matrem meam et sic sequar te dixitque ei uade et

19:21 And he returned back from him, and took a yoke reuertere quod enim meum erat feci tibi 19:21 reuersus au-

of oxen, and slew them, and boiled their flesh with the instruments of the oxen, and gave unto the people, and they did eat. Then he arose, and went after Elijah, and ministered unto him.

tem ab eo tulit par boum et mactauit illud et in aratro boum coxit carnes et dedit populo et comederunt consurgensque abiit et secutus est Heliam et ministrabat ei

20:1 porro Benadad rex Syriae congregauit omnem exerci- 20:1 And Benhadad the king of Syria gathered all his host

together: and there were thirty and two kings with him, and horses, and chariots; and he went up and besieged

tum suum et triginta et duos reges secum et equos et currus Samaria, and warred against it. et ascendens pugnabat contra Samariam et obsidebat eam 20:2 mittensque nuntios ad Ahab regem Israhel in ciuitatem 20:2 And he sent messengers to Ahab king of Israel into

the city, and said unto him, Thus saith Benhadad,

20:3 ait haec dicit Benadad argentum tuum et aurum tuum 20:3 Thy silver and thy gold is mine; thy wives also and

thy children, even the goodliest, are mine.

meum est et uxores tuae et filii tui optimi mei sunt 20:4 res- 20:4 And the king of Israel answered and said, My lord,

O king, according to thy saying, I am thine, and all that I have.

ponditque rex Israhel iuxta uerbum tuum domine mi rex tuus sum ego et omnia mea 20:5 reuertentesque nuntii dixerunt 20:5 And the messengers came again, and said, Thus

speaketh Benhadad, saying, Although I have sent unto thee, saying, Thou shalt deliver me thy silver, and thy gold,

haec dicit Benadad qui misit nos ad te argentum tuum et au- and thy wives, and thy children; rum tuum et uxores tuas et filios tuos dabis mihi 20:6 cras 20:6 Yet I will send my servants unto thee to morrow about

this time, and they shall search thine house, and the houses of thy servants; and it shall be, that whatsoever is pleasant in thine eyes, they shall put it in their hand, and take it away.

igitur hac eadem hora mittam seruos meos ad te et scrutabuntur domum tuam et domum seruorum tuorum et omne quod eis placuerit ponent in manibus suis et auferent 20:7 uoca- 20:7 Then the king of Israel called all the elders of the

land, and said, Mark, I pray you, and see how this man seeketh mischief: for he sent unto me for my wives, and for my children, and for my silver, and for my gold; and I denied him not.

uit autem rex Israhel omnes seniores terrae et ait animaduertite et uidete quoniam insidietur nobis misit enim ad me pro uxoribus meis et filiis et pro argento et auro et non abnui 20:8 dixeruntque omnes maiores natu et uniuersus populus 20:8 And all the elders and all the people said unto him,

Hearken not unto him, nor consent.

ad eum non audias neque adquiescas illi 20:9 respondit ita- 20:9 Wherefore he said unto the messengers of Benhadad,

Tell my lord the king, All that thou didst send for to thy servant at the first I will do: but this thing I may not do. And the messengers departed, and brought him word again.

que nuntiis Benadad dicite domino meo regi omnia propter quae misisti ad me seruum tuum initio faciam hanc autem rem facere non possum 20:10 reuersique nuntii rettulerunt 20:10 And Benhadad sent unto him, and said, The gods

do so unto me, and more also, if the dust of Samaria shall suffice for handfuls for all the people that follow me.

ei qui remisit et ait haec faciant mihi dii et haec addant si suffecerit puluis Samariae pugillis omnis populi qui sequitur me 20:11 et respondens rex Israhel ait dicite ei ne glo- 20:11 And the king of Israel answered and said, Tell him,

Let not him that girdeth on his harness boast himself as he that putteth it off.

rietur accinctus aeque ut discinctus 20:12 factum est autem

20:12 And it came to pass, when Ben-hadad heard this message, as he was drinking, he and the kings in the pavilions, that he said unto his servants, Set yourselves in array. And they set themselves in array against the city.

cum audisset uerbum istud bibebat ipse et reges in umbraculis et ait seruis suis circumdate ciuitatem et circumdederunt eam 20:13 et ecce propheta unus accedens ad Ahab regem 20:13 And, behold, there came a prophet unto Ahab king

of Israel, saying, Thus saith the LORD, Hast thou seen all this great multitude? behold, I will deliver it into thine

Israhel ait haec dicit Dominus certe uidisti omnem multitu- hand this day; and thou shalt know that I am the LORD. dinem hanc nimiam ecce ego tradam eam in manu tua hodie ut scias quia ego sum Dominus 20:14 et ait Ahab per 20:14 And Ahab said, By whom? And he said, Thus saith

the LORD, Even by the young men of the princes of the provinces. Then he said, Who shall order the battle? And

quem dixitque ei haec dicit Dominus per pedisequos princi- he answered, Thou.

Biblia Sacra Vulgata

pum prouinciarum et ait quis incipiet proeliari et ille dixit tu

20:15 Then he numbered the young men of the princes of 20:15 recensuit ergo pueros principum prouinciarum et rep-

the provinces, and they were two hundred and thirty two: and after them he numbered all the people, even all the

children of Israel, being seven thousand. perit numerum ducentorum triginta duum et post eos recen-20:16 And they went out at noon. But Benhadad was suit populum omnes filios Israhel septem milia 20:16 et eg-

drinking himself drunk in the pavilions, he and the kings, the thirty and two kings that helped him.

ressi sunt meridie Benadad autem bibebat temulentus in umbraculo suo et reges triginta duo cum eo qui ad auxilium eius

20:17 And the young men of the princes of the provinces uenerant 20:17 egressi sunt autem pueri principum prouin-

went out first; and Benhadad sent out, and they told him, saying, There are men come out of Samaria.

ciarum in prima fronte misit itaque Benadad qui nuntiaue-

20:18 And he said, Whether they be come out for peace, runt ei dicentes uiri egressi sunt de Samaria 20:18 at ille

take them alive; or whether they be come out for war, take them alive.

siue ait pro pace ueniunt adprehendite eos uiuos siue ut proe-

20:19 So these young men of the princes of the provinces lientur uiuos eos capite 20:19 egressi sunt ergo pueri princi-

came out of the city, and the army which followed them.

20:20 And they slew every one his man: and the Syrians pum prouinciarum ac reliquus exercitus sequebatur 20:20 et

fled; and Israel pursued them: and Benhadad the king of Syria escaped on an horse with the horsemen.

percussit unusquisque uirum qui contra se uenerat fugeruntque Syri et persecutus est eos Israhel fugit quoque Benadad

20:21 And the king of Israel went out, and smote the rex Syriae in equo cum equitibus 20:21 necnon et egres-

horses and chariots, and slew the Syrians with a great slaughter.

sus rex Israhel percussit equos et currus et percussit Syriam

20:22 And the prophet came to the king of Israel, and said plaga magna 20:22 accedens autem propheta ad regem Isra-

unto him, Go, strengthen thyself, and mark, and see what thou doest: for at the return of the year the king of Syria

will come up against thee. hel dixit ei uade et confortare et scito et uide quid facias se-20:23 And the servants of the king of Syria said unto quenti enim anno rex Syriae ascendet contra te 20:23 serui

him, Their gods are gods of the hills; therefore they were stronger than we; but let us fight against them in the plain,

and surely we shall be stronger than they. uero regis Syriae dixerunt ei dii montium sunt dii eorum ideo

superauerunt nos sed melius est ut pugnemus contra eos in

20:24 And do this thing, Take the kings away, every man campestribus et obtinebimus eos 20:24 tu ergo uerbum hoc

out of his place, and put captains in their rooms:

fac amoue reges singulos ab exercitu suo et pone principes

20:25 And number thee an army, like the army that thou pro eis 20:25 et instaura numerum militum qui ceciderunt

hast lost, horse for horse, and chariot for chariot: and we will fight against them in the plain, and surely we shall be stronger than they. And he hearkened unto their voice, and did so.

de tuis et equos secundum equos pristinos et currus secundum currus quos ante habuisti et pugnabimus contra eos in campestribus et uidebis quod obtinebimus eos credidit con-

20:26 And it came to pass at the return of the year, that silio eorum et fecit ita 20:26 igitur postquam annus transie-

Benhadad numbered the Syrians, and went up to Aphek, to fight against Israel.

rat recensuit Benadad Syros et ascendit in Afec ut pugnaret

20:27 And the children of Israel were numbered, and were contra Israhel 20:27 porro filii Israhel recensiti sunt et ac-

all present, and went against them: and the children of Israel pitched before them like two little flocks of kids;

but the Syrians filled the country. ceptis cibariis profecti ex aduerso castraque metati contra

eos quasi duo parui greges caprarum Syri autem repleuerunt

20:28 And there came a man of God, and spake unto the terram 20:28 et accedens unus uir Dei dixit ad regem Is-

king of Israel, and said, Thus saith the LORD, Because the Syrians have said, The LORD is God of the hills, but he is not God of the valleys, therefore will I deliver all this great multitude into thine hand, and ye shall know that I am the LORD.

rahel haec dicit Dominus quia dixerunt Syri deus montium est Dominus et non est deus uallium dabo omnem multitudinem grandem hanc in manu tua et scietis quia ego Dominus 20:29 dirigebant septem diebus ex aduerso hii atque illi acies 20:29 And they pitched one over against the other seven

days. And so it was, that in the seventh day the battle was joined: and the children of Israel slew of the Syrians an

septima autem die commissum est bellum percusseruntque hundred thousand footmen in one day. filii Israhel de Syris centum milia peditum in die una 20:30 20:30 But the rest fled to Aphek, into the city; and there a

wall fell upon twenty and seven thousand of the men that were left. And Benhadad fled, and came into the city, into

fugerunt autem qui remanserant in Afec in ciuitatem et ceci- an inner chamber. dit murus super uiginti septem milia hominum qui remanserant porro Benadad fugiens ingressus est ciuitatem in cubiculum quod erat intra cubiculum 20:31 dixeruntque ei serui 20:31 And his servants said unto him, Behold now, we

have heard that the kings of the house of Israel are merciful kings: let us, I pray thee, put sackcloth on our loins, and ropes upon our heads, and go out to the king of Israel: peradventure he will save thy life.

sui ecce audiuimus quod reges domus Israhel clementes sint ponamus itaque saccos in lumbis nostris et funiculos in capitibus nostris et egrediamur ad regem Israhel forsitan saluabit animas nostras 20:32 accinxerunt saccis lumbos suos et po- 20:32 So they girded sackcloth on their loins, and put

ropes on their heads, and came to the king of Israel, and said, Thy servant Benhadad saith, I pray thee, let me live.

suerunt funes in capitibus ueneruntque ad regem Israhel et And he said, Is he yet alive? he is my brother. dixerunt seruus tuus Benadad dicit uiuat oro te anima mea et ille ait si adhuc uiuit frater meus est 20:33 quod acceperunt 20:33 Now the men did diligently observe whether any

thing would come from him, and did hastily catch it: and they said, Thy brother Benhadad. Then he said, Go ye, bring him. Then Benhadad came forth to him; and he caused him to come up into the chariot.

uiri pro omine et festinantes rapuerunt uerbum ex ore eius atque dixerunt frater tuus Benadad et dixit eis ite et adducite eum egressus est ergo ad eum Benadad et leuauit eum in currum suum 20:34 qui dixit ei ciuitates quas tulit pater 20:34 And Ben-hadad said unto him, The cities, which my

father took from thy father, I will restore; and thou shalt make streets for thee in Damascus, as my father made in Samaria. Then said Ahab, I will send thee away with this covenant. So he made a covenant with him, and sent him away.

meus a patre tuo reddam et plateas fac tibi in Damasco sicut fecit pater meus in Samaria et ego foederatus recedam a te pepigit ergo foedus et dimisit eum 20:35 tunc uir quidam 20:35 And a certain man of the sons of the prophets said

unto his neighbour in the word of the LORD, Smite me, I pray thee. And the man refused to smite him.

de filiis prophetarum dixit ad socium suum in sermone Domini percute me at ille noluit percutere 20:36 cui ait quia 20:36 Then said he unto him, Because thou hast not

obeyed the voice of the LORD, behold, as soon as thou art departed from me, a lion shall slay thee. And as soon as he was departed from him, a lion found him, and slew him.

noluisti audire uocem Domini ecce recedes a me et percutiet te leo cumque paululum recessisset ab eo inuenit eum leo atque percussit 20:37 sed et alterum conueniens uirum dixit 20:37 Then he found another man, and said, Smite me, I

pray thee. And the man smote him, so that in smiting he wounded him.

ad eum percute me qui percussit eum et uulnerauit 20:38

20:38 So the prophet departed, and waited for the king by the way, and disguised himself with ashes upon his face.

abiit ergo propheta et occurrit regi in uia et mutauit aspersione pulueris os et oculos suos 20:39 cumque rex transiret 20:39 And as the king passed by, he cried unto the king:

and he said, Thy servant went out into the midst of the battle; and, behold, a man turned aside, and brought a man unto me, and said, Keep this man: if by any means he be missing, then shall thy life be for his life, or else thou shalt pay a talent of silver.

clamauit ad regem et ait seruus tuus egressus est ad proe-

Biblia Sacra Vulgata

liandum comminus cumque fugisset uir unus adduxit eum quidam ad me et ait custodi uirum istum qui si lapsus fuerit erit anima tua pro anima eius aut talentum argenti adpendes

20:40 And as thy servant was busy here and there, he was 20:40 dum autem ego turbatus huc illucque me uerterem su-

gone. And the king of Israel said unto him, So shall thy judgment be; thyself hast decided it.

bito non conparuit et ait rex Israhel ad eum hoc est iudicium

20:41 And he hasted, and took the ashes away from his tuum quod ipse decreuisti 20:41 at ille statim abstersit pul-

face; and the king of Israel discerned him that he was of the prophets.

uerem de facie sua et cognouit eum rex Israhel quod esset de

20:42 And he said unto him, Thus saith the LORD, Be- prophetis 20:42 qui ait ad eum haec dicit Dominus quia di-

cause thou hast let go out of thy hand a man whom I appointed to utter destruction, therefore thy life shall go for

his life, and thy people for his people. misisti uirum dignum morte de manu tua erit anima tua pro 20:43 And the king of Israel went to his house heavy and anima eius et populus tuus pro populo eius 20:43 reuersus

displeased, and came to Samaria.

est igitur rex Israhel in domum suam audire contemnens et furibundus uenit Samariam

21:1 And it came to pass after these things, that Naboth 21:1 post uerba autem haec uinea erat Naboth Hiezraheli-

the Jezreelite had a vineyard, which was in Jezreel, hard by the palace of Ahab king of Samaria.

tae qui erat in Hiezrahel iuxta palatium Ahab regis Samariae

21:2 And Ahab spake unto Naboth, saying, Give me thy 21:2 locutus est ergo Ahab ad Naboth dicens da mihi uineam

vineyard, that I may have it for a garden of herbs, because it is near unto my house: and I will give thee for it a better vineyard than it; or, if it seem good to thee, I will give thee the worth of it in money.

tuam ut faciam mihi hortum holerum quia uicina est et prope domum meam daboque tibi pro ea uineam meliorem aut si

21:3 And Naboth said to Ahab, The LORD forbid it me, tibi commodius putas argenti pretium quanto digna est 21:3

that I should give the inheritance of my fathers unto thee.

cui respondit Naboth propitius mihi sit Dominus ne dem he-

21:4 And Ahab came into his house heavy and displeased reditatem patrum meorum tibi 21:4 uenit ergo Ahab in do-

because of the word which Naboth the Jezreelite had spoken to him: for he had said, I will not give thee the inheritance of my fathers. And he laid him down upon his bed, and turned away his face, and would eat no bread.

mum suam indignans et frendens super uerbo quod locutus fuerat ad eum Naboth Hiezrahelites dicens non do tibi hereditatem patrum meorum et proiciens se in lectulum suum

21:5 But Jezebel his wife came to him, and said unto him, auertit faciem ad parietem et non comedit panem 21:5 in-

Why is thy spirit so sad, that thou eatest no bread?

gressa est autem ad eum Hiezabel uxor sua dixitque ei quid est hoc unde anima tua contristata est et quare non comedis

21:6 And he said unto her, Because I spake unto Naboth panem 21:6 qui respondit ei locutus sum Naboth Hiezra-

the Jezreelite, and said unto him, Give me thy vineyard for money; or else, if it please thee, I will give thee another vineyard for it: and he answered, I will not give thee my vineyard.

helitae et dixi ei da mihi uineam tuam accepta pecunia aut si tibi placet dabo tibi uineam pro ea et ille ait non do tibi

21:7 And Jezebel his wife said unto him, Dost thou now uineam meam 21:7 dixit ergo ad eum Hiezabel uxor eius govern the kingdom of Israel? arise, and eat bread, and let thine heart be merry: I will give thee the vineyard of

Naboth the Jezreelite. grandis auctoritatis es et bene regis regnum Israhel surge et

comede panem et aequo esto animo ego dabo tibi uineam

21:8 So she wrote letters in Ahab’s name, and sealed them Naboth Hiezrahelitae 21:8 scripsit itaque litteras ex nomine

with his seal, and sent the letters unto the elders and to the nobles that were in his city, dwelling with Naboth.

Ahab et signauit eas anulo eius et misit ad maiores natu et ad optimates qui erant in ciuitate eius et habitabant cum Naboth 21:9 litterarum autem erat ista sententia praedicate ieiunium 21:9 And she wrote in the letters, saying, Proclaim a fast,

and set Naboth on high among the people:

et sedere facite Naboth inter primos populi 21:10 et sub- 21:10 And set two men, sons of Belial, before him, to bear

witness against him, saying, Thou didst blaspheme God and the king. And then carry him out, and stone him, that

mittite duos uiros filios Belial contra eum et falsum testi- he may die. monium dicant benedixit Deum et regem et educite eum et lapidate sicque moriatur 21:11 fecerunt ergo ciues eius ma- 21:11 And the men of his city, even the elders and the

nobles who were the inhabitants in his city, did as Jezebel had sent unto them, and as it was written in the letters

iores natu et optimates qui habitabant cum eo in urbe sic- which she had sent unto them. ut praeceperat eis Hiezabel et sicut scriptum erat in litteris quas miserat ad eos 21:12 praedicauerunt ieiunium et se- 21:12 They proclaimed a fast, and set Naboth on high

among the people.

dere fecerunt Naboth inter primos populi 21:13 et adductis 21:13 And there came in two men, children of Belial, and

sat before him: and the men of Belial witnessed against him, even against Naboth, in the presence of the people, saying, Naboth did blaspheme God and the king. Then they carried him forth out of the city, and stoned him with stones, that he died.

duobus uiris filiis diaboli fecerunt eos sedere contra eum at illi scilicet ut uiri diabolici dixerunt contra eum testimonium coram multitudine benedixit Naboth Deo et regi quam ob rem eduxerunt eum extra ciuitatem et lapidibus interfecerunt 21:14 miseruntque ad Hiezabel dicentes lapidatus est 21:14 Then they sent to Jezebel, saying, Naboth is stoned,

and is dead.

Naboth et mortuus est 21:15 factum est autem cum audisset 21:15 And it came to pass, when Jezebel heard that Naboth

was stoned, and was dead, that Jezebel said to Ahab, Arise, take possession of the vineyard of Naboth the Jezreelite, which he refused to give thee for money: for Naboth is not alive, but dead.

Hiezabel lapidatum Naboth et mortuum locuta est ad Ahab surge posside uineam Naboth Hiezrahelitae qui noluit tibi adquiescere et dare eam accepta pecunia non enim uiuit Naboth sed mortuus est 21:16 quod cum audisset Ahab mor- 21:16 And it came to pass, when Ahab heard that Naboth

was dead, that Ahab rose up to go down to the vineyard of Naboth the Jezreelite, to take possession of it.

tuum uidelicet Naboth surrexit et descendebat in uineam Naboth Hiezrahelitae ut possideret eam 21:17 factus est igitur 21:17 And the word of the LORD came to Elijah the Tish-

bite, saying,

sermo Domini ad Heliam Thesbiten dicens 21:18 surge et 21:18 Arise, go down to meet Ahab king of Israel, which

is in Samaria: behold, he is in the vineyard of Naboth, whither he is gone down to possess it.

descende in occursum Ahab regis Israhel qui est in Samaria ecce ad uineam Naboth descendit ut possideat eam 21:19 et 21:19 And thou shalt speak unto him, saying, Thus saith

the LORD, Hast thou killed, and also taken possession? And thou shalt speak unto him, saying, Thus saith the LORD, In the place where dogs licked the blood of Naboth shall dogs lick thy blood, even thine.

loqueris ad eum dicens haec dicit Dominus occidisti insuper et possedisti et post haec addes haec dicit Dominus in loco hoc in quo linxerunt canes sanguinem Naboth lambent tuum quoque sanguinem 21:20 et ait Ahab ad Heliam num inue- 21:20 And Ahab said to Elijah, Hast thou found me, O

mine enemy? And he answered, I have found thee: because thou hast sold thyself to work evil in the sight of the

nisti me inimice mee qui dixit inueni eo quod uenundatus LORD. sis ut faceres malum in conspectu Domini 21:21 ecce ego 21:21 Behold, I will bring evil upon thee, and will take

away thy posterity, and will cut off from Ahab him that pisseth against the wall, and him that is shut up and left in

inducam super te malum et demetam posteriora tua et inter- Israel,

Biblia Sacra Vulgata

ficiam de Ahab mingentem ad parietem et clausum et ulti-

21:22 And will make thine house like the house of Jer- mum in Israhel 21:22 et dabo domum tuam sicut domum oboam the son of Nebat, and like the house of Baasha the son of Ahijah, for the provocation wherewith thou hast

provoked me to anger, and made Israel to sin. Hieroboam filii Nabath et sicut domum Baasa filii Ahia quia

egisti ut me ad iracundiam prouocares et peccare fecisti Is-

21:23 And of Jezebel also spake the LORD, saying, The rahel 21:23 sed et de Hiezabel locutus est Dominus dicens

dogs shall eat Jezebel by the wall of Jezreel. 21:24 Him that dieth of Ahab in the city the dogs shall eat; canes comedent Hiezabel in agro Hiezrahel 21:24 si morand him that dieth in the field shall the fowls of the air eat.

tuus fuerit Ahab in ciuitate comedent eum canes si autem

21:25 But there was none like unto Ahab, which did sell mortuus fuerit in agro comedent eum uolucres caeli 21:25

himself to work wickedness in the sight of the LORD, whom Jezebel his wife stirred up.

igitur non fuit alter talis ut Ahab qui uenundatus est ut faceret malum in conspectu Domini concitauit enim eum Hie-

21:26 And he did very abominably in following idols, ac- zabel uxor sua 21:26 et abominabilis effectus est in tantum

cording to all things as did the Amorites, whom the LORD cast out before the children of Israel.

ut sequeretur idola quae fecerant Amorrei quos consumpsit

21:27 And it came to pass, when Ahab heard those words, Dominus a facie filiorum Israhel 21:27 itaque cum audis-

that he rent his clothes, and put sackcloth upon his flesh, and fasted, and lay in sackcloth, and went softly.

set Ahab sermones istos scidit uestem suam et operuit cilicio carnem suam ieiunauitque et dormiuit in sacco et ambu-

21:28 And the word of the LORD came to Elijah the Tish- labat dimisso capite 21:28 factus est autem sermo Domini

bite, saying,

21:29 Seest thou how Ahab humbleth himself before me? ad Heliam Thesbiten dicens 21:29 nonne uidisti humilia-

because he humbleth himself before me, I will not bring the evil in his days: but in his son’s days will I bring the

evil upon his house. tum Ahab coram me quia igitur humiliatus est mei causa non inducam malum in diebus eius sed in diebus filii sui inferam malum domui eius

22:1 And they continued three years without war between 22:1 transierunt igitur tres anni absque bello inter Syriam

Syria and Israel.

22:2 And it came to pass in the third year, that Jehoshaphat et Israhel 22:2 in anno autem tertio descendit Iosaphat rex

the king of Judah came down to the king of Israel.

22:3 And the king of Israel said unto his servants, Know Iuda ad regem Israhel 22:3 dixitque rex Israhel ad seruos

ye that Ramoth in Gilead is ours, and we be still, and take it not out of the hand of the king of Syria?

suos ignoratis quod nostra sit Ramoth Galaad et neglegimus

22:4 And he said unto Jehoshaphat, Wilt thou go with me tollere eam de manu regis Syriae 22:4 et ait ad Iosaphat

to battle to Ramothgilead? And Jehoshaphat said to the king of Israel, I am as thou art, my people as thy people,

my horses as thy horses. ueniesne mecum ad proeliandum in Ramoth Galaad 22:5

22:5 And Jehoshaphat said unto the king of Israel, Enquire, I pray thee, at the word of the LORD to day.

dixitque Iosaphat ad regem Israhel sicut ego sum ita et tu populus meus et populus tuus unum sunt et equites mei et equites tui dixitque Iosaphat ad regem Israhel quaere oro

22:6 Then the king of Israel gathered the prophets to- te hodie sermonem Domini 22:6 congregauit ergo rex Is-

gether, about four hundred men, and said unto them, Shall I go against Ramothgilead to battle, or shall I forbear? And they said, Go up; for the LORD shall deliver it into the hand of the king.

rahel prophetas quadringentos circiter uiros et ait ad eos ire debeo in Ramoth Galaad ad bellandum an quiescere qui re-

22:7 And Jehoshaphat said, Is there not here a prophet of sponderunt ascende et dabit Dominus in manu regis 22:7 the LORD besides, that we might enquire of him?

dixit autem Iosaphat non est hic propheta Domini quispiam ut interrogemus per eum 22:8 et ait rex Israhel ad Iosaphat 22:8 And the king of Israel said unto Jehoshaphat, There

is yet one man, Micaiah the son of Imlah, by whom we may enquire of the LORD: but I hate him; for he doth not prophesy good concerning me, but evil. And Jehoshaphat said, Let not the king say so.

remansit uir unus per quem possimus interrogare Dominum sed ego odi eum quia non prophetat mihi bonum sed malum Micheas filius Hiemla cui Iosaphat ait ne loquaris ita rex 22:9 uocauit ergo rex Israhel eunuchum quendam et dixit ei 22:9 Then the king of Israel called an officer, and said,

Hasten hither Micaiah the son of Imlah.

festina adducere Micheam filium Hiemla 22:10 rex autem 22:10 And the king of Israel and Jehoshaphat the king of

Judah sat each on his throne, having put on their robes, in a void place in the entrance of the gate of Samaria; and all

Israhel et Iosaphat rex Iuda sedebat unusquisque in solio suo the prophets prophesied before them. uestiti cultu regio in area iuxta ostium portae Samariae et uniuersi prophetae prophetabant in conspectu eorum 22:11 22:11 And Zedekiah the son of Chenaanah made him

horns of iron: and he said, Thus saith the LORD, With these shalt thou push the Syrians, until thou have con-

fecit quoque sibi Sedecias filius Chanaan cornua ferrea et sumed them. ait haec dicit Dominus his uentilabis Syriam donec deleas eam 22:12 omnesque prophetae similiter prophetabant di- 22:12 And all the prophets prophesied so, saying, Go up

to Ramothgilead, and prosper: for the LORD shall deliver it into the king’s hand.

centes ascende in Ramoth Galaad et uade prospere et tradet Dominus in manu regis 22:13 nuntius uero qui ierat ut 22:13 And the messenger that was gone to call Micaiah

spake unto him, saying, Behold now, the words of the prophets declare good unto the king with one mouth: let thy word, I pray thee, be like the word of one of them, and speak that which is good.

uocaret Micheam locutus est ad eum dicens ecce sermones prophetarum ore uno bona regi praedicant sit ergo et sermo tuus similis eorum et loquere bona 22:14 cui Micheas ait 22:14 And Micaiah said, As the LORD liveth, what the

LORD saith unto me, that will I speak.

uiuit Dominus quia quodcumque dixerit mihi Dominus hoc loquar 22:15 uenit itaque ad regem et ait illi rex Michea ire 22:15 So he came to the king. And the king said unto him,

Micaiah, shall we go against Ramothgilead to battle, or shall we forbear? And he answered him, Go, and prosper:

debemus in Ramoth Galaad ad proeliandum an cessare cui for the LORD shall deliver it into the hand of the king. ille respondit ascende et uade prospere et tradet Dominus in manu regis 22:16 dixit autem rex ad eum iterum atque 22:16 And the king said unto him, How many times shall I

adjure thee that thou tell me nothing but that which is true in the name of the LORD?

iterum adiuro te ut non loquaris mihi nisi quod uerum est

in nomine Domini 22:17 et ille ait uidi cunctum Israhel di- 22:17 And he said, I saw all Israel scattered upon the hills, as sheep that have not a shepherd: and the LORD said, These have no master: let them return every man to his

spersum in montibus quasi oues non habentes pastorem et house in peace. ait Dominus non habent dominum isti reuertatur unusquis-

que in domum suam in pace 22:18 dixit ergo rex Israhel ad 22:18 And the king of Israel said unto Jehoshaphat, Did I not tell thee that he would prophesy no good concerning me, but evil?

Iosaphat numquid non dixi tibi quia non prophetat mihi bonum sed semper malum 22:19 ille uero addens ait propterea 22:19 And he said, Hear thou therefore the word of the

LORD: I saw the LORD sitting on his throne, and all the host of heaven standing by him on his right hand and on

audi sermonem Domini uidi Dominum sedentem super so- his left. lium suum et omnem exercitum caeli adsistentem ei a dextris et a sinistris 22:20 et ait Dominus quis decipiet Ahab regem 22:20 And the LORD said, Who shall persuade Ahab, that

he may go up and fall at Ramothgilead? And one said on this manner, and another said on that manner.

Biblia Sacra Vulgata

Israhel ut ascendat et cadat in Ramoth Galaad et dixit unus

22:21 And there came forth a spirit, and stood before the uerba huiuscemodi et alius aliter 22:21 egressus est autem

LORD, and said, I will persuade him.

spiritus et stetit coram Domino et ait ego decipiam illum cui

22:22 And the LORD said unto him, Wherewith? And he locutus est Dominus in quo 22:22 et ille ait egrediar et ero

said, I will go forth, and I will be a lying spirit in the mouth of all his prophets. And he said, Thou shalt persude him,

and prevail also: go forth, and do so. spiritus mendax in ore omnium prophetarum eius et dixit 22:23 Now therefore, behold, the LORD hath put a lying Dominus decipies et praeualebis egredere et fac ita 22:23

spirit in the mouth of all these thy prophets, and the LORD hath spoken evil concerning thee.

nunc igitur ecce dedit Dominus spiritum mendacii in ore omnium prophetarum tuorum qui hic sunt et Dominus locu-

22:24 But Zedekiah the son of Chenaanah went near, and tus est contra te malum 22:24 accessit autem Sedecias filius

smote Micaiah on the cheek, and said, Which way went the Spirit of the LORD from me to speak unto thee?

Chanaan et percussit Micheam in maxillam et dixit mene

22:25 And Micaiah said, Behold, thou shalt see in that day, ergo dimisit spiritus Domini et locutus est tibi 22:25 et ait

when thou shalt go into an inner chamber to hide thyself.

Micheas uisurus es in die illa quando ingredieris cubiculum

22:26 And the king of Israel said, Take Micaiah, and carry intra cubiculum ut abscondaris 22:26 et ait rex Israhel tol-

him back unto Amon the governor of the city, and to Joash the king’s son;

lite Micheam et maneat apud Amon principem ciuitatis et

22:27 And say, Thus saith the king, Put this fellow in the apud Ioas filium Ammelech 22:27 et dicite eis haec dicit prison, and feed him with bread of affliction and with water of affliction, until I come in peace.

rex mittite uirum istum in carcerem et sustentate eum pane

22:28 And Micaiah said, If thou return at all in peace, the tribulationis et aqua angustiae donec reuertar in pace 22:28

LORD hath not spoken by me. And he said, Hearken, O people, every one of you.

dixitque Micheas si reuersus fueris in pace non est locutus

22:29 So the king of Israel and Jehoshaphat the king of Dominus in me et ait audite populi omnes 22:29 ascendit

Judah went up to Ramothgilead.

itaque rex Israhel et Iosaphat rex Iuda in Ramoth Galaad

22:30 And the king of Israel said unto Jehoshaphat, I will 22:30 dixitque rex Israhel ad Iosaphat sume arma et ingre-

disguise myself, and enter into the battle; but put thou on thy robes. And the king of Israel disguised himself, and

went into the battle. dere proelium et induere uestibus tuis porro rex Israhel muta-22:31 But the king of Syria commanded his thirty and two uit habitum et ingressus est bellum 22:31 rex autem Syriae

captains that had rule over his chariots, saying, Fight neither with small nor great, save only with the king of Israel.

praeceperat principibus curruum triginta duobus dicens non pugnabitis contra minorem et maiorem quempiam nisi con-

22:32 And it came to pass, when the captains of the char- tra regem Israhel solum 22:32 cum ergo uidissent principes

iots saw Jehoshaphat, that they said, Surely it is the king of Israel. And they turned aside to fight against him: and

Jehoshaphat cried out. curruum Iosaphat suspicati sunt quod ipse esset rex Israhel

et impetu facto pugnabant contra eum et exclamauit Iosaphat

22:33 And it came to pass, when the captains of the char- 22:33 intellexeruntque principes curruum quod non esset rex

iots perceived that it was not the king of Israel, that they turned back from pursuing him.

Israhel et cessauerunt ab eo 22:34 unus autem quidam te-

22:34 And a certain man drew a bow at a venture, and smote the king of Israel between the joints of the harness: wherefore he said unto the driver of his chariot, Turn thine hand, and carry me out of the host; for I am wounded.

tendit arcum in incertum sagittam dirigens et casu percussit regem Israhel inter pulmonem et stomachum at ille dixit aurigae suo uerte manum tuam et eice me de exercitu quia grauiter uulneratus sum 22:35 commissum est ergo proe- 22:35 And the battle increased that day: and the king was

stayed up in his chariot against the Syrians, and died at even: and the blood ran out of the wound into the midst of

lium in die illa et rex Israhel stabat in curru suo contra Sy- the chariot. ros et mortuus est uesperi fluebat autem sanguis plagae in sinum currus 22:36 et praeco personuit in uniuerso exer- 22:36 And there went a proclamation throughout the host

about the going down of the sun, saying, Every man to his city, and every man to his own country.

citu antequam sol occumberet dicens unusquisque reuertatur in ciuitatem et in terram suam 22:37 mortuus est autem 22:37 So the king died, and was brought to Samaria; and

they buried the king in Samaria.

rex et perlatus est Samariam sepelieruntque regem in Samaria 22:38 et lauerunt currum in piscina Samariae et linxe- 22:38 And one washed the chariot in the pool of Samaria;

and the dogs licked up his blood; and they washed his armour; according unto the word of the LORD which he

runt canes sanguinem eius et habenas lauerunt iuxta uerbum spake. Domini quod locutus fuerat 22:39 reliqua uero sermonum 22:39 Now the rest of the acts of Ahab, and all that he did,

and the ivory house which he made, and all the cities that he built, are they not written in the book of the chronicles

Ahab et uniuersa quae fecit et domus eburneae quam aedifi- of the kings of Israel? cauit cunctarumque urbium quas extruxit nonne scripta sunt haec in libro uerborum dierum regum Israhel 22:40 dormi- 22:40 So Ahab slept with his fathers; and Ahaziah his son

reigned in his stead.

uit ergo Ahab cum patribus suis et regnauit Ohozias filius eius pro eo 22:41 Iosaphat filius Asa regnare coeperat super 22:41 And Jehoshaphat the son of Asa began to reign over

Judah in the fourth year of Ahab king of Israel.

Iudam anno quarto Ahab regis Israhel 22:42 triginta quin- 22:42 Jehoshaphat was thirty and five years old when he

began to reign; and he reigned twenty and five years in Jerusalem. And his mother’s name was Azubah the daugh-

que annorum erat cum regnare coepisset et uiginti et quin- ter of Shilhi. que annos regnauit in Hierusalem nomen matris eius Azuba filia Salai 22:43 et ambulauit in omni uia Asa patris sui et 22:43 And he walked in all the ways of Asa his father; he

turned not aside from it, doing that which was right in the eyes of the LORD:

non declinauit ex ea fecitque quod rectum est in conspectu Domini 22:44 uerumtamen excelsa non abstulit adhuc enim 22:44 nevertheless the high places were not taken away;

for the people offered and burnt incense yet in the high places.

populus sacrificabat et adolebat incensum in excelsis 22:45

22:45 And Jehoshaphat made peace with the king of Israel.

pacemque habuit Iosaphat cum rege Israhel 22:46 reliqua 22:46 Now the rest of the acts of Jehoshaphat, and his

might that he shewed, and how he warred, are they not written in the book of the chronicles of the kings of Ju-

autem uerborum Iosaphat et opera eius quae gessit et pro- dah? elia nonne haec scripta sunt in libro uerborum dierum regum Iuda 22:47 sed et reliquias effeminatorum qui reman- 22:47 And the remnant of the sodomites, which remained

in the days of his father Asa, he took out of the land.

serant in diebus Asa patris eius abstulit de terra 22:48 nec 22:48 There was then no king in Edom: a deputy was king. erat tunc rex constitutus in Edom 22:49 rex uero Iosaphat 22:49 Jehoshaphat made ships of Tharshish to go to Ophir

for gold: but they went not; for the ships were broken at Eziongeber.

fecerat classes in mari quae nauigarent in Ophir propter aurum et ire non potuerunt quia confractae sunt in Asiongaber 22:50 tunc ait Ohozias filius Ahab ad Iosaphat uadant serui 22:50 Then said Ahaziah the son of Ahab unto Je-

hoshaphat, Let my servants go with thy servants in the ships. But Jehoshaphat would not.

mei cum seruis tuis in nauibus et noluit Iosaphat 22:51 dor-

22:51 And Jehoshaphat slept with his fathers, and was buried with his fathers in the city of David his father: and Jehoram his son reigned in his stead.

miuitque cum patribus suis et sepultus est cum eis in ciuitate

Biblia Sacra Vulgata

22:52 Ahaziah the son of Ahab began to reign over Israel Dauid patris sui regnauitque Ioram filius eius pro eo 22:52

in Samaria the seventeenth year of Jehoshaphat king of Judah, and reigned two years over Israel.

Ohozias autem filius Ahab regnare coeperat super Israhel in Samaria anno septimodecimo Iosaphat regis Iuda regna-

22:53 And he did evil in the sight of the LORD, and uitque super Israhel duobus annis 22:53 et fecit malum in

walked in the way of his father, and in the way of his mother, and in the way of Jeroboam the son of Nebat, who

made Israel to sin: conspectu Domini et ambulauit in uia patris sui et matris

suae et in uia Hieroboam filii Nabath qui peccare fecit Isra-

22:54 For he served Baal, and worshipped him, and pro- hel 22:54 seruiuit quoque Baal et adorauit eum et inritauit

voked to anger the LORD God of Israel, according to all that his father had done.

Dominum Deum Israhel iuxta omnia quae fecerat pater eius

Liber II Regum

1:1 praeuaricatus est autem Moab in Israhel postquam 1:1 Then Moab rebelled against Israel after the death of

Ahab.

mortuus est Ahab 1:2 ceciditque Ohozias per cancellos ce- 1:2 And Ahaziah fell down through a lattice in his upper

chamber that was in Samaria, and was sick: and he sent messengers, and said unto them, Go, enquire of Baalzebub

naculi sui quod habebat in Samaria et aegrotauit misitque the god of Ekron whether I shall recover of this disease. nuntios dicens ad eos ite consulite Beelzebub deum Accaron utrum uiuere queam de infirmitate mea hac 1:3 angelus au- 1:3 But the angel of the LORD said to Elijah the Tish-

bite, Arise, go up to meet the messengers of the king of Samaria, and say unto them, Is it not because there is not a God in Israel, that ye go to enquire of Baalzebub the god of Ekron?

tem Domini locutus est ad Heliam Thesbiten surge ascende in occursum nuntiorum regis Samariae et dices ad eos numquid non est Deus in Israhel ut eatis ad consulendum Beelzebub deum Accaron 1:4 quam ob rem haec dicit Dominus 1:4 Now therefore thus saith the LORD, Thou shalt not

come down from that bed on which thou art gone up, but shalt surely die. And Elijah departed.

de lectulo super quem ascendisti non descendes sed morte morieris et abiit Helias 1:5 reuersique sunt nuntii ad Oho- 1:5 And when the messengers turned back unto him, he

said unto them, Why are ye now turned back?

ziam qui dixit eis quare reuersi estis 1:6 at illi responderunt 1:6 And they said unto him, There came a man up to meet

us, and said unto us, Go, turn again unto the king that sent you, and say unto him, Thus saith the LORD, Is it not because there is not a God in Israel, that thou sendest to enquire of Baalzebub the god of Ekron? therefore thou shalt not come down from that bed on which thou art gone up, but shalt surely die.

ei uir occurrit nobis et dixit ad nos ite reuertimini ad regem qui misit uos et dicetis ei haec dicit Dominus numquid quia non erat Deus in Israhel mittis ut consulatur Beelzebub deus Accaron idcirco de lectulo super quem ascendisti non descendes sed morte morieris 1:7 qui dixit eis cuius figurae 1:7 And he said unto them, What manner of man was he

which came up to meet you, and told you these words?

et habitu est uir qui occurrit uobis et locutus est uerba haec 1:8 at illi dixerunt uir pilosus et zona pellicia accinctis reni- 1:8 And they answered him, He was an hairy man, and girt

with a girdle of leather about his loins. And he said, It is Elijah the Tishbite.

bus qui ait Helias Thesbites est 1:9 misitque ad eum quin-

1:9 Then the king sent unto him a captain of fifty with his fifty. And he went up to him: and, behold, he sat on the top of an hill. And he spake unto him, Thou man of God, the king hath said, Come down.

quagenarium principem et quinquaginta qui erant sub eo qui ascendit ad eum sedentique in uertice montis ait homo Dei rex praecepit ut descendas 1:10 respondensque Helias dixit 1:10 And Elijah answered and said to the captain of fifty,

If I be a man of God, then let fire come down from heaven, and consume thee and thy fifty. And there came down fire

quinquagenario si homo Dei sum descendat ignis e caelo et from heaven, and consumed him and his fifty. deuoret te et quinquaginta tuos descendit itaque ignis e caelo et deuorauit eum et quinquaginta qui erant cum eo 1:11 rur- 1:11 Again also he sent unto him another captain of fifty

with his fifty. And he answered and said unto him, O man of God, thus hath the king said, Come down quickly.

sum misit ad eum principem quinquagenarium alterum et quinquaginta cum eo qui locutus est illi homo Dei haec dicit rex festina descende 1:12 respondens Helias ait si homo 1:12 And Elijah answered and said unto them, If I be a

man of God, let fire come down from heaven, and consume thee and thy fifty. And the fire of God came down

Dei ego sum descendat ignis e caelo et deuoret te et quinqua- from heaven, and consumed him and his fifty. ginta tuos descendit ergo ignis Dei e caelo et deuorauit illum et quinquaginta eius 1:13 iterum misit principem quinqua- 1:13 And he sent again a captain of the third fifty with his

fifty. And the third captain of fifty went up, and came and fell on his knees before Elijah, and besought him, and said unto him, O man of God, I pray thee, let my life, and the life of these fifty thy servants, be precious in thy sight.

genarium tertium et quinquaginta qui erant cum eo qui cum

Biblia Sacra Vulgata

uenisset curuauit genua contra Heliam et precatus est eum et ait homo Dei noli despicere animam meam et animam ser-

1:14 Behold, there came fire down from heaven, and burnt uorum tuorum qui mecum sunt 1:14 ecce descendit ignis de

up the two captains of the former fifties with their fifties: therefore let my life now be precious in thy sight.

caelo et deuorauit duos principes quinquagenarios primos et quinquagenos qui cum eis erant sed nunc obsecro ut mise-

1:15 And the angel of the LORD said unto Elijah, Go rearis animae meae 1:15 locutus est autem angelus Domini

down with him: be not afraid of him. And he arose, and went down with him unto the king.

ad Heliam dicens descende cum eo ne timeas surrexit igitur

1:16 And he said unto him, Thus saith the LORD, Foras- et descendit cum eo ad regem 1:16 et locutus est ei haec di-

much as thou hast sent messengers to enquire of Baalzebub the god of Ekron, is it not because there is no God in Israel to enquire of his word? therefore thou shalt not come down off that bed on which thou art gone up, but shalt surely die.

cit Dominus quia misisti nuntios ad consulendum Beelzebub deum Accaron quasi non esset Deus in Israhel a quo possis interrogare sermonem ideo de lectulo super quem ascendisti

1:17 So he died according to the word of the LORD which non descendes sed morte morieris 1:17 mortuus est ergo

Elijah had spoken. And Jehoram reigned in his stead in the second year of Jehoram the son of Jehoshaphat king of

Judah; because he had no son. iuxta sermonem Domini quem locutus est Helias et regnauit

Ioram frater eius pro eo anno secundo Ioram filii Iosaphat re-

1:18 Now the rest of the acts of Ahaziah which he did, are gis Iudae non enim habebat filium 1:18 reliqua autem uer-

they not written in the book of the chronicles of the kings of Israel?

borum Ohoziae quae operatus est nonne haec scripta sunt in libro sermonum dierum regum Israhel

2:1 And it came to pass, when the LORD would take up 2:1 factum est autem cum leuare uellet Dominus Heliam

Elijah into heaven by a whirlwind, that Elijah went with Elisha from Gilgal.

per turbinem in caelum ibant Helias et Heliseus de Galga-

2:2 And Elijah said unto Elisha, Tarry here, I pray thee; for lis 2:2 dixitque Helias ad Heliseum sede hic quia Domithe LORD hath sent me to Bethel. And Elisha said unto him, As the LORD liveth, and as thy soul liveth, I will not

leave thee. So they went down to Bethel. nus misit me usque Bethel cui ait Heliseus uiuit Dominus

et uiuit anima tua quia non derelinquam te cumque descen-

2:3 And the sons of the prophets that were at Bethel came dissent Bethel 2:3 egressi sunt filii prophetarum qui erant

forth to Elisha, and said unto him, Knowest thou that the LORD will take away thy master from thy head to day?

And he said, Yea, I know it; hold ye your peace. Bethel ad Heliseum et dixerunt ei numquid nosti quia hodie

Dominus tollat dominum tuum a te qui respondit et ego noui

2:4 And Elijah said unto him, Elisha, tarry here, I pray silete 2:4 dixit autem Helias ad Heliseum sede hic quia Do-

thee; for the LORD hath sent me to Jericho. And he said, As the LORD liveth, and as thy soul liveth, I will not leave

thee. So they came to Jericho. minus misit me in Hiericho et ille ait uiuit Dominus et uiuit

anima tua quia non derelinquam te cumque uenissent Hie-

2:5 And the sons of the prophets that were at Jericho came richum 2:5 accesserunt filii prophetarum qui erant in Hie-

to Elisha, and said unto him, Knowest thou that the LORD will take away thy master from thy head to day? And he

answered, Yea, I know it; hold ye your peace. richo ad Heliseum et dixerunt ei numquid nosti quia hodie

Dominus tollet dominum tuum a te et ait et ego noui silete

2:6 And Elijah said unto him, Tarry, I pray thee, here; for 2:6 dixit autem ei Helias sede hic quia Dominus misit me the LORD hath sent me to Jordan. And he said, As the LORD liveth, and as thy soul liveth, I will not leave thee.

And they two went on. ad Iordanem qui ait uiuit Dominus et uiuit anima tua quia

Liber II Regum

non derelinquam te ierunt igitur ambo pariter 2:7 et quin- 2:7 And fifty men of the sons of the prophets went, and

stood to view afar off: and they two stood by Jordan.

quaginta uiri de filiis prophetarum secuti sunt qui et steterunt e contra longe illi autem ambo stabant super Iordanem 2:8 tulitque Helias pallium suum et inuoluit illud et percus- 2:8 And Elijah took his mantle, and wrapped it together,

and smote the waters, and they were divided hither and thither, so that they two went over on dry ground.

sit aquas quae diuisae sunt in utramque partem et transierunt ambo per siccum 2:9 cumque transissent Helias dixit 2:9 And it came to pass, when they were gone over, that

Elijah said unto Elisha, Ask what I shall do for thee, before I be taken away from thee. And Elisha said, I pray thee,

ad Heliseum postula quod uis ut faciam tibi antequam tol- let a double portion of thy spirit be upon me. lar a te dixitque Heliseus obsecro ut fiat duplex spiritus tuus in me 2:10 qui respondit rem difficilem postulasti attamen 2:10 And he said, Thou hast asked a hard thing: neverthe-

less, if thou see me when I am taken from thee, it shall be so unto thee; but if not, it shall not be so.

si uideris me quando tollor a te erit quod petisti si autem non uideris non erit 2:11 cumque pergerent et incedentes 2:11 And it came to pass, as they still went on, and talked,

that, behold, there appeared a chariot of fire, and horses of fire, and parted them both asunder; and Elijah went up by

sermocinarentur ecce currus igneus et equi ignei diuiserunt a whirlwind into heaven.

utrumque et ascendit Helias per turbinem in caelum 2:12 2:12 And Elisha saw it, and he cried, My father, my father, the chariot of Israel, and the horsemen thereof. And he saw him no more: and he took hold of his own clothes,

Heliseus autem uidebat et clamabat pater mi pater mi currus and rent them in two pieces. Israhel et auriga eius et non uidit eum amplius adprehenditque uestimenta sua et scidit illa in duas partes 2:13 et 2:13 He took up also the mantle of Elijah that fell from

him, and went back, and stood by the bank of Jordan;

leuauit pallium Heliae quod ceciderat ei reuersusque stetit super ripam Iordanis 2:14 et pallio Heliae quod ceciderat 2:14 And he took the mantle of Elijah that fell from him,

and smote the waters, and said, Where is the LORD God of Elijah? and when he also had smitten the waters, they

ei percussit aquas et dixit ubi est Deus Heliae etiam nunc parted hither and thither: and Elisha went over. percussitque aquas et diuisae sunt huc atque illuc et transiit Heliseus 2:15 uidentes autem filii prophetarum qui erant 2:15 And when the sons of the prophets which were to

view at Jericho saw him, they said, The spirit of Elijah doth rest on Elisha. And they came to meet him, and

in Hiericho de contra dixerunt requieuit spiritus Heliae su- bowed themselves to the ground before him. per Heliseum et uenientes in occursum eius adorauerunt eum proni in terram 2:16 dixeruntque illi ecce cum seruis tuis 2:16 And they said unto him, Behold now, there be with

thy servants fifty strong men; let them go, we pray thee, and seek thy master: lest peradventure the Spirit of the LORD hath taken him up, and cast him upon some mountain, or into some valley. And he said, Ye shall not send.

sunt quinquaginta uiri fortes qui possint ire et quaerere dominum tuum ne forte tulerit eum spiritus Domini et proiecerit in uno montium aut in una uallium qui ait nolite mittere 2:17 coegeruntque eum donec adquiesceret et diceret 2:17 And when they urged him till he was ashamed, he

said, Send. They sent therefore fifty men; and they sought three days, but found him not.

mittite et miserunt quinquaginta uiros qui cum quaesissent tribus diebus non inuenerunt 2:18 et reuersi sunt ad eum 2:18 And when they came again to him, (for he tarried at

Jericho,) he said unto them, Did I not say unto you, Go not?

at ille habitabat in Hiericho dixitque eis numquid non dixi uobis nolite ire 2:19 dixerunt quoque uiri ciuitatis ad Heli- 2:19 And the men of the city said unto Elisha, Behold, I

pray thee, the situation of this city is pleasant, as my lord seeth: but the water is naught, and the ground barren.

seum ecce habitatio ciuitatis huius optima est sicut tu ipse

Biblia Sacra Vulgata

domine perspicis sed aquae pessimae sunt et terra sterilis

2:20 And he said, Bring me a new cruse, and put salt 2:20 at ille ait adferte mihi uas nouum et mittite in illud sal

therein. And they brought it to him.

2:21 And he went forth unto the spring of the waters, and qui cum adtulissent 2:21 egressus ad fontem aquarum misit

cast the salt in there, and said, Thus saith the LORD, I have healed these waters; there shall not be from thence

any more death or barren land. in eum sal et ait haec dicit Dominus sanaui aquas has et non 2:22 So the waters were healed unto this day, according to erit ultra in eis mors neque sterilitas 2:22 sanatae sunt ergo

the saying of Elisha which he spake.

aquae usque ad diem hanc iuxta uerbum Helisei quod locu-

2:23 And he went up from thence unto Bethel: and as he tus est 2:23 ascendit autem inde Bethel cumque ascenderet

was going up by the way, there came forth little children out of the city, and mocked him, and said unto him, Go up,

thou bald head; go up, thou bald head. per uiam pueri parui egressi sunt de ciuitate et inludebant ei 2:24 And he turned back, and looked on them, and cursed dicentes ascende calue ascende calue 2:24 qui cum se res-

them in the name of the LORD. And there came forth two she bears out of the wood, and tare forty and two children

of them. pexisset uidit eos et maledixit eis in nomine Domini egressi-

que sunt duo ursi de saltu et lacerauerunt ex eis quadraginta

2:25 And he went from thence to mount Carmel, and from duos pueros 2:25 abiit autem inde in montem Carmeli et

thence he returned to Samaria.

inde reuersus est Samariam

3:1 Now Jehoram the son of Ahab began to reign over 3:1 Ioram uero filius Ahab regnauit super Israhel in Sa-

Israel in Samaria the eighteenth year of Jehoshaphat king of Judah, and reigned twelve years.

maria anno octauodecimo Iosaphat regis Iudae regnauitque

3:2 And he wrought evil in the sight of the LORD; but not duodecim annis 3:2 et fecit malum coram Domino sed non like his father, and like his mother: for he put away the image of Baal that his father had made.

sicut pater suus et mater tulit enim statuas Baal quas fecerat

3:3 Nevertheless he cleaved unto the sins of Jeroboam the pater eius 3:3 uerumtamen in peccatis Hieroboam filii Nason of Nebat, which made Israel to sin; he departed not therefrom.

bath qui peccare fecit Israhel adhesit nec recessit ab eis 3:4

3:4 And Mesha king of Moab was a sheepmaster, and rendered unto the king of Israel an hundred thousand lambs, and an hundred thousand rams, with the wool.

porro Mesa rex Moab nutriebat pecora multa et soluebat regi Israhel centum milia agnorum et centum milia arietum cum

3:5 But it came to pass, when Ahab was dead, that the king uelleribus suis 3:5 cumque mortuus fuisset Ahab praeuari-

of Moab rebelled against the king of Israel.

3:6 And king Jehoram went out of Samaria the same time, catus est foedus quod habebat cum rege Israhel 3:6 egres-

and numbered all Israel.

sus est igitur rex Ioram in die illa de Samaria et recensuit uni-

3:7 And he went and sent to Jehoshaphat the king of Judah, uersum Israhel 3:7 misitque ad Iosaphat regem Iuda dicens

saying, The king of Moab hath rebelled against me: wilt thou go with me against Moab to battle? And he said, I will go up: I am as thou art, my people as thy people, and my horses as thy horses.

rex Moab recessit a me ueni mecum contra Moab ad proelium qui respondit ascendam qui meus est tuus est populus

3:8 And he said, Which way shall we go up? And he an- meus populus tuus equi mei equi tui 3:8 dixitque per quam

swered, The way through the wilderness of Edom.

uiam ascendemus at ille respondit per desertum Idumeae

3:9 So the king of Israel went, and the king of Judah, and 3:9 perrexerunt igitur rex Israhel et rex Iuda et rex Edom

the king of Edom: and they fetched a compass of seven days’ journey: and there was no water for the host, and for

the cattle that followed them. et circumierunt per uiam septem dierum nec erat aqua ex-3:10 And the king of Israel said, Alas! that the LORD hath ercitui et iumentis quae sequebantur eos 3:10 dixitque rex

called these three kings together, to deliver them into the hand of Moab!

Israhel eheu eheu eheu congregauit nos Dominus tres reges

Liber II Regum

ut traderet in manu Moab 3:11 et ait Iosaphat estne hic pro- 3:11 But Jehoshaphat said, Is there not here a prophet of

the LORD, that we may enquire of the LORD by him? And one of the king of Israel’s servants answered and said, Here is Elisha the son of Shaphat, which poured water on the hands of Elijah.

pheta Domini ut deprecemur Dominum per eum et respondit unus de seruis regis Israhel est hic Heliseus filius Saphat qui fundebat aquam super manus Heliae 3:12 et ait Iosaphat est 3:12 And Jehoshaphat said, The word of the LORD is with

him. So the king of Israel and Jehoshaphat and the king of Edom went down to him.

apud eum sermo Domini descenditque ad eum rex Israhel et Iosaphat et rex Edom 3:13 dixit autem Heliseus ad regem 3:13 And Elisha said unto the king of Israel, What have

I to do with thee? get thee to the prophets of thy father, and to the prophets of thy mother. And the king of Israel said unto him, Nay: for the LORD hath called these three kings together, to deliver them into the hand of Moab.

Israhel quid mihi et tibi est uade ad prophetas patris tui et matris tuae et ait illi rex Israhel quare congregauit Dominus tres reges hos ut traderet eos in manu Moab 3:14 dixit au- 3:14 And Elisha said, As the LORD of hosts liveth, before

whom I stand, surely, were it not that I regard the presence of Jehoshaphat the king of Judah, I would not look toward

tem Heliseus uiuit Dominus exercituum in cuius conspectu thee, nor see thee. sto quod si non uultum Iosaphat regis Iudae erubescerem ne adtendissem quidem te nec respexissem 3:15 nunc autem 3:15 But now bring me a minstrel. And it came to pass,

when the minstrel played, that the hand of the LORD came upon him.

adducite mihi psalten cumque caneret psaltes facta est super eum manus Domini et ait 3:16 haec dicit Dominus facite 3:16 And he said, Thus saith the LORD, Make this valley

full of ditches.

alueum torrentis huius fossas et fossas 3:17 haec enim di- 3:17 For thus saith the LORD, Ye shall not see wind, nei-

ther shall ye see rain; yet that valley shall be filled with water, that ye may drink, both ye, and your cattle, and

cit Dominus non uidebitis uentum neque pluuiam et alueus your beasts. iste replebitur aquis et bibetis uos et familiae uestrae et iu-

menta uestra 3:18 parumque hoc est in conspectu Domini 3:18 And this is but a light thing in the sight of the LORD:

he will deliver the Moabites also into your hand.

insuper tradet etiam Moab in manu uestra 3:19 et percu- 3:19 And ye shall smite every fenced city, and every

choice city, and shall fell every good tree, and stop all wells of water, and mar every good piece of land with

tietis omnem ciuitatem munitam et omnem urbem electam stones. et uniuersum lignum fructiferum succidetis cunctosque fontes aquarum obturabitis et omnem agrum egregium operietis lapidibus 3:20 factum est igitur mane quando sacrificium 3:20 And it came to pass in the morning, when the meat

offering was offered, that, behold, there came water by the way of Edom, and the country was filled with water.

offerri solet et ecce aquae ueniebant per uiam Edom et repleta est terra aquis 3:21 uniuersi autem Moabitae audientes 3:21 And when all the Moabites heard that the kings were

come up to fight against them, they gathered all that were able to put on armour, and upward, and stood in the border.

quod ascendissent reges ut pugnarent aduersum eos conuocauerunt omnes qui accincti erant balteo desuper et steterunt in terminis 3:22 primoque mane surgentes et orto iam sole 3:22 And they rose up early in the morning, and the sun

shone upon the water, and the Moabites saw the water on the other side as red as blood:

ex aduerso aquarum uiderunt Moabitae contra aquas rubras quasi sanguinem 3:23 dixeruntque sanguis est gladii pug- 3:23 And they said, This is blood: the kings are surely

slain, and they have smitten one another: now therefore, Moab, to the spoil.

nauerunt reges contra se et caesi sunt mutuo nunc perge ad praedam Moab 3:24 perrexeruntque in castra Israhel porro 3:24 And when they came to the camp of Israel, the Is-

raelites rose up and smote the Moabites, so that they fled before them: but they went forward smiting the Moabites,

consurgens Israhel percussit Moab at illi fugerunt coram eis even in their country.

Biblia Sacra Vulgata

3:25 And they beat down the cities, and on every good uenerunt igitur qui uicerant et percusserunt Moab 3:25 et

piece of land cast every man his stone, and filled it; and they stopped all the wells of water, and felled all the good trees: only in Kirharaseth left they the stones thereof; howbeit the slingers went about it, and smote it.

ciuitates destruxerunt et omnem agrum optimum mittentes singuli lapides repleuerunt et uniuersos fontes aquarum obturauerunt et omnia ligna fructifera succiderunt ita ut muri tantum fictiles remanerent et circumdata est ciuitas a fundi-

3:26 And when the king of Moab saw that the battle was balariis et magna ex parte percussa 3:26 quod cum uidisset

too sore for him, he took with him seven hundred men that drew swords, to break through even unto the king of

Edom: but they could not. rex Moab praeualuisse scilicet hostes tulit secum septingen-

tos uiros educentes gladium ut inrumperet ad regem Edom et

3:27 Then he took his eldest son that should have reigned non potuerunt 3:27 arripiensque filium suum primogenitum

in his stead, and offered him for a burnt offering upon the wall. And there was great indignation against Israel: and

they departed from him, and returned to their own land. qui regnaturus erat pro eo obtulit holocaustum super murum et facta est indignatio magna in Israhel statimque recesserunt ab eo et reuersi sunt in terram suam

4:1 Now there cried a certain woman of the wives of the 4:1 mulier autem quaedam de uxoribus prophetarum cla-

sons of the prophets unto Elisha, saying, Thy servant my husband is dead; and thou knowest that thy servant did fear the LORD: and the creditor is come to take unto him my two sons to be bondmen.

mabat ad Heliseum dicens seruus tuus uir meus mortuus est et tu nosti quia seruus tuus fuit timens Dominum et ecce creditor uenit ut tollat duos filios meos ad seruiendum sibi

4:2 And Elisha said unto her, What shall I do for thee? 4:2 cui dixit Heliseus quid uis ut faciam tibi dic mihi quid

tell me, what hast thou in the house? And she said, Thine handmaid hath not any thing in the house, save a pot of

oil. habes in domo tua at illa respondit non habeo ancilla tua

4:3 Then he said, Go, borrow thee vessels abroad of all thy quicquam in domo mea nisi parum olei quo unguear 4:3 neighbours, even empty vessels; borrow not a few.

cui ait uade pete mutuo ab omnibus uicinis tuis uasa uacua

4:4 And when thou art come in, thou shalt shut the door non pauca 4:4 et ingredere et claude ostium cum intrinse-

upon thee and upon thy sons, and shalt pour out into all those vessels, and thou shalt set aside that which is full.

cus fueris tu et filii tui et mitte inde in omnia uasa haec et

4:5 So she went from him, and shut the door upon her and cum plena fuerint tolles 4:5 iuit itaque mulier et clusit os-

upon her sons, who brought the vessels to her; and she poured out.

tium super se et super filios suos illi offerebant uasa et illa

4:6 And it came to pass, when the vessels were full, that infundebat 4:6 cumque plena fuissent uasa dixit ad filium

she said unto her son, Bring me yet a vessel. And he said unto her, There is not a vessel more. And the oil stayed.

suum adfer mihi adhuc uas et ille respondit non habeo ste-

4:7 Then she came and told the man of God. And he said, titque oleum 4:7 uenit autem illa et indicauit homini Dei

Go, sell the oil, and pay thy debt, and live thou and thy children of the rest.

et ille uade inquit uende oleum et redde creditori tuo tu au-

4:8 And it fell on a day, that Elisha passed to Shunem, tem et filii tui uiuite de reliquo 4:8 facta est autem quae-

where was a great woman; and she constrained him to eat bread. And so it was, that as oft as he passed by, he turned

in thither to eat bread. dam dies et transiebat Heliseus per Sunam erat autem ibi mulier magna quae tenuit eum ut comederet panem cumque frequenter inde transiret deuertebat ad eam ut comede-

4:9 And she said unto her husband, Behold now, I per- ret panem 4:9 quae dixit ad uirum suum animaduerto quod

ceive that this is an holy man of God, which passeth by us continually.

Liber II Regum

uir Dei sanctus est iste qui transit per nos frequenter 4:10 4:10 Let us make a little chamber, I pray thee, on the wall;

and let us set for him there a bed, and a table, and a stool, and a candlestick: and it shall be, when he cometh to us,

faciamus ergo cenaculum paruum et ponamus ei in eo lec- that he shall turn in thither. tulum et mensam et sellam et candelabrum ut cum uenerit ad nos maneat ibi 4:11 facta est igitur dies quaedam et ue- 4:11 And it fell on a day, that he came thither, and he

turned into the chamber, and lay there.

niens deuertit in cenaculum et requieuit ibi 4:12 dixitque ad 4:12 And he said to Gehazi his servant, Call this Shunam-

mite. And when he had called her, she stood before him.

Giezi puerum suum uoca Sunamitin istam qui cum uocasset eam et illa stetisset coram eo 4:13 dixit ad puerum loquere 4:13 And he said unto him, Say now unto her, Behold,

thou hast been careful for us with all this care; what is to be done for thee? wouldest thou be spoken for to the king, or to the captain of the host? And she answered, I dwell among mine own people.

ad eam ecce sedule in omnibus ministrasti nobis quid uis ut faciam tibi numquid habes negotium et uis ut loquar regi siue principi militiae quae respondit in medio populi mei habito 4:14 et ait quid ergo uult ut faciam ei dixitque Giezi 4:14 And he said, What then is to be done for her? And

Gehazi answered, Verily she hath no child, and her husband is old.

ne quaeras filium enim non habet et uir eius senex est 4:15

4:15 And he said, Call her. And when he had called her, she stood in the door.

praecepit itaque ut uocaret eam quae cum uocata fuisset et stetisset ad ostium 4:16 dixit ad eam in tempore isto et in 4:16 And he said, About this season, according to the time

of life, thou shalt embrace a son. And she said, Nay, my lord, thou man of God, do not lie unto thine handmaid.

hac eadem hora si uita comes fuerit habebis in utero filium at illa respondit noli quaeso domine mi uir Dei noli mentiri ancillae tuae 4:17 et concepit mulier et peperit filium in 4:17 And the woman conceived, and bare a son at that

season that Elisha had said unto her, according to the time of life.

tempore et in hora eadem quam dixerat Heliseus 4:18 cre-

4:18 And when the child was grown, it fell on a day, that he went out to his father to the reapers.

uit autem puer et cum esset quaedam dies et egressus isset ad patrem suum ad messores 4:19 ait patri suo caput meum 4:19 And he said unto his father, My head, my head. And

he said to a lad, Carry him to his mother.

caput meum at ille dixit puero tolle et duc eum ad matrem suam 4:20 qui cum tulisset et adduxisset eum ad matrem 4:20 And when he had taken him, and brought him to his

mother, he sat on her knees till noon, and then died.

suam posuit eum illa super genua sua usque ad meridiem et mortuus est 4:21 ascendit autem et conlocauit eum super 4:21 And she went up, and laid him on the bed of the man

of God, and shut the door upon him, and went out.

lectulum hominis Dei et clusit ostium et egressa 4:22 uoca- 4:22 And she called unto her husband, and said, Send me,

I pray thee, one of the young men, and one of the asses, that I may run to the man of God, and come again.

uit uirum suum et ait mitte mecum obsecro unum de pueris et asinam ut excurram usque ad hominem Dei et reuertar 4:23 4:23 And he said, Wherefore wilt thou go to him to day?

it is neither new moon, nor sabbath. And she said, It shall be well.

qui ait illi quam ob causam uadis ad eum hodie non sunt kalendae neque sabbatum quae respondit uale 4:24 strauitque 4:24 Then she saddled an ass, and said to her servant,

Drive, and go forward; slack not thy riding for me, except I bid thee.

asinam et praecepit puero mina et propera ne mihi moram facias in eundo et hoc age quod praecipio tibi 4:25 profecta 4:25 So she went and came unto the man of God to mount

Carmel. And it came to pass, when the man of God saw her afar off, that he said to Gehazi his servant, Behold,

est igitur et uenit ad uirum Dei in montem Carmeli cumque yonder is that Shunammite: uidisset eam uir Dei de contra ait ad Giezi puerum suum ecce

Biblia Sacra Vulgata

4:26 Run now, I pray thee, to meet her, and say unto her, Sunamitis illa 4:26 uade ergo in occursum eius et dic ei re-

Is it well with thee? is it well with thy husband? is it well with the child? And she answered, It is well:

ctene agitur circa te et circa uirum tuum et circa filium tuum

4:27 And when she came to the man of God to the hill, quae respondit recte 4:27 cumque uenisset ad uirum Dei in

she caught him by the feet: but Gehazi came near to thrust her away. And the man of God said, Let her alone; for her soul is vexed within her: and the LORD hath hid it from me, and hath not told me.

monte adprehendit pedes eius et accessit Giezi ut amoueret eam et ait homo Dei dimitte illam anima enim eius in amari-

4:28 Then she said, Did I desire a son of my lord? did I tudine est et Dominus celauit me et non indicauit mihi 4:28

not say, Do not deceive me?

quae dixit illi numquid petiui filium a domino meo numquid

4:29 Then he said to Gehazi, Gird up thy loins, and take non dixi tibi ne inludas me 4:29 et ille ait ad Giezi accinge

my staff in thine hand, and go thy way: if thou meet any man, salute him not; and if any salute thee, answer him

not again: and lay my staff upon the face of the child. lumbos tuos et tolle baculum meum in manu tua et uade si occurrerit tibi homo non salutes eum et si salutauerit te quispiam non respondeas illi et pones baculum meum super fa-

4:30 And the mother of the child said, As the LORD liveth, ciem pueri 4:30 porro mater pueri ait uiuit Dominus et uiuit

and as thy soul liveth, I will not leave thee. And he arose, and followed her.

anima tua non dimittam te surrexit ergo et secutus est eam

4:31 And Gehazi passed on before them, and laid the staff 4:31 Giezi autem praecesserat eos et posuerat baculum su-

upon the face of the child; but there was neither voice, nor hearing. Wherefore he went again to meet him, and told

him, saying, The child is not awaked. per faciem pueri et non erat uox neque sensus reuersusque

est in occursum eius et nuntiauit ei dicens non surrexit puer

4:32 And when Elisha was come into the house, behold, 4:32 ingressus est ergo Heliseus domum et ecce puer mor-

the child was dead, and laid upon his bed.

4:33 He went in therefore, and shut the door upon them tuus iacebat in lectulo eius 4:33 ingressusque clusit ostium

twain, and prayed unto the LORD.

4:34 And he went up, and lay upon the child, and put his super se et puerum et orauit ad Dominum 4:34 et ascendit

mouth upon his mouth, and his eyes upon his eyes, and his hands upon his hands: and stretched himself upon the

child; and the flesh of the child waxed warm. et incubuit super puerum posuitque os suum super os eius et oculos suos super oculos eius et manus suas super manus eius et incuruauit se super eum et calefacta est caro pueri

4:35 Then he returned, and walked in the house to and 4:35 at ille reuersus deambulauit in domo semel huc et illuc

fro; and went up, and stretched himself upon him: and the child sneezed seven times, and the child opened his eyes.

et ascendit et incubuit super eum et oscitauit puer septies

4:36 And he called Gehazi, and said, Call this Shunam- aperuitque oculos 4:36 et ille uocauit Giezi et dixit ei uoca

mite. So he called her. And when she was come in unto him, he said, Take up thy son.

Sunamitin hanc quae uocata ingressa est ad eum qui ait tolle

4:37 Then she went in, and fell at his feet, and bowed her- filium tuum 4:37 uenit illa et corruit ad pedes eius et adora-

self to the ground, and took up her son, and went out.

4:38 And Elisha came again to Gilgal: and there was a uit super terram tulitque filium suum et egressa est 4:38 et

dearth in the land; and the sons of the prophets were sitting before him: and he said unto his servant, Set on the great

pot, and seethe pottage for the sons of the prophets. Heliseus reuersus est in Galgala erat autem fames in terra et filii prophetarum habitabant coram eo dixitque uni de pueris suis pone ollam grandem et coque pulmentum filiis pro-

4:39 And one went out into the field to gather herbs, and phetarum 4:39 et egressus est unus in agrum ut colligeret

found a wild vine, and gathered thereof wild gourds his lap full, and came and shred them into the pot of pottage:

for they knew them not. herbas agrestes inuenitque quasi uitem siluestrem et colle-

Liber II Regum

git ex ea colocyntidas agri et impleuit pallium suum et reuersus concidit in ollam pulmenti nesciebat enim quid esset 4:40 infuderunt ergo sociis ut comederent cumque gustassent 4:40 So they poured out for the men to eat. And it came

to pass, as they were eating of the pottage, that they cried out, and said, O thou man of God, there is death in the pot.

de coctione exclamauerunt dicentes mors in olla uir Dei et And they could not eat thereof. non potuerunt comedere 4:41 at ille adferte inquit farinam 4:41 But he said, Then bring meal. And he cast it into the

pot; and he said, Pour out for the people, that they may eat. And there was no harm in the pot.

et misit in ollam et ait infunde turbae et comedat et non fuit amplius quicquam amaritudinis in olla 4:42 uir autem qui- 4:42 And there came a man from Baalshalisha, and

brought the man of God bread of the firstfruits, twenty loaves of barley, and full ears of corn in the husk thereof.

dam uenit de Balsalisa deferens uiro Dei panes primitiarum And he said, Give unto the people, that they may eat. et uiginti panes hordiacios et frumentum nouum in pera sua at ille dixit da populo ut comedat 4:43 responditque ei mi- 4:43 And his servitor said, What, should I set this before

an hundred men? He said again, Give the people, that they may eat: for thus saith the LORD, They shall eat, and shall

nister eius quantum est hoc ut adponam coram centum uiris leave thereof. rursum ille da ait populo ut comedat haec enim dicit Dominus comedent et supererit 4:44 posuit itaque coram eis qui 4:44 So he set it before them, and they did eat, and left

thereof, according to the word of the LORD.

comederunt et superfuit iuxta uerbum Domini

5:1 Naaman princeps militiae regis Syriae erat uir mag- 5:1 Now Naaman, captain of the host of the king of Syria,

was a great man with his master, and honourable, because by him the LORD had given deliverance unto Syria: he

nus apud dominum suum et honoratus per illum enim dedit was also a mighty man in valour, but he was a leper. Dominus salutem Syriae erat autem uir fortis et diues sed leprosus 5:2 porro de Syria egressi fuerant latrunculi et ca- 5:2 And the Syrians had gone out by companies, and had

brought away captive out of the land of Israel a little maid; and she waited on Naaman’s wife.

ptiuam duxerant de terra Israhel puellam paruulam quae erat in obsequio uxoris Naaman 5:3 quae ait ad dominam suam 5:3 And she said unto her mistress, Would God my lord

were with the prophet that is in Samaria! for he would recover him of his leprosy.

utinam fuisset dominus meus ad prophetam qui est in Samaria profecto curasset eum a lepra quam habet 5:4 ingressus 5:4 And one went in, and told his lord, saying, Thus and

thus said the maid that is of the land of Israel.

est itaque Naaman ad dominum suum et nuntiauit ei dicens sic et sic locuta est puella de terra Israhel 5:5 dixitque ei 5:5 And the king of Syria said, Go to, go, and I will send

a letter unto the king of Israel. And he departed, and took with him ten talents of silver, and six thousand pieces of

rex Syriae uade et mittam litteras ad regem Israhel qui cum gold, and ten changes of raiment. profectus esset et tulisset secum decem talenta argenti et sex milia aureos et decem mutatoria uestimentorum 5:6 detulit 5:6 And he brought the letter to the king of Israel, say-

ing, Now when this letter is come unto thee, behold, I have therewith sent Naaman my servant to thee, that thou

litteras ad regem Israhel in haec uerba cum acceperis epis- mayest recover him of his leprosy. tulam hanc scito quod miserim ad te Naaman seruum meum ut cures eum a lepra sua 5:7 cumque legisset rex Israhel 5:7 And it came to pass, when the king of Israel had read

the letter, that he rent his clothes, and said, Am I God, to kill and to make alive, that this man doth send unto me to recover a man of his leprosy? wherefore consider, I pray you, and see how he seeketh a quarrel against me.

litteras scidit uestimenta sua et ait numquid Deus sum ut occidere possim et uiuificare quia iste misit ad me ut curem hominem a lepra sua animaduertite et uidete quod occasio-

Biblia Sacra Vulgata

5:8 And it was so, when Elisha the man of God had heard nes quaerat aduersum me 5:8 quod cum audisset Heliseus

that the king of Israel had rent his clothes, that he sent to the king, saying, Wherefore hast thou rent thy clothes? let him come now to me, and he shall know that there is a prophet in Israel.

uir Dei scidisse uidelicet regem Israhel uestimenta sua misit ad eum dicens quare scidisti uestimenta tua ueniat ad me et

5:9 So Naaman came with his horses and with his chariot, sciat esse prophetam in Israhel 5:9 uenit ergo Naaman cum

and stood at the door of the house of Elisha.

5:10 And Elisha sent a messenger unto him, saying, Go equis et curribus et stetit ad ostium domus Helisei 5:10 mi-

and wash in Jordan seven times, and thy flesh shall come again to thee, and thou shalt be clean.

sitque ad eum Heliseus nuntium dicens uade et lauare septies in Iordane et recipiet sanitatem caro tua atque mundaberis

5:11 But Naaman was wroth, and went away, and said, 5:11 iratus Naaman recedebat dicens putabam quod egrede-

Behold, I thought, He will surely come out to me, and stand, and call on the name of the LORD his God, and

strike his hand over the place, and recover the leper. retur ad me et stans inuocaret nomen Domini Dei sui et tan-

5:12 Are not Abana and Pharpar, rivers of Damascus, bet- geret manu sua locum leprae et curaret me 5:12 numquid ter than all the waters of Israel? may I not wash in them, and be clean? So he turned and went away in a rage.

non meliores sunt Abana et Pharphar fluuii Damasci omnibus aquis Israhel ut lauer in eis et munder cum ergo uertis-

5:13 And his servants came near, and spake unto him, and set se et abiret indignans 5:13 accesserunt ad eum serui sui

said, My father, if the prophet had bid thee do some great thing, wouldest thou not have done it? how much rather

then, when he saith to thee, Wash, and be clean? et locuti sunt ei pater si rem grandem dixisset tibi propheta

certe facere debueras quanto magis quia nunc dixit tibi la-

5:14 Then went he down, and dipped himself seven times uare et mundaberis 5:14 descendit et lauit in Iordane sep-

in Jordan, according to the saying of the man of God: and his flesh came again like unto the flesh of a little child, and

he was clean. ties iuxta sermonem uiri Dei et restituta est caro eius sicut 5:15 And he returned to the man of God, he and all his caro pueri paruuli et mundatus est 5:15 reuersusque ad ui-

company, and came, and stood before him: and he said, Behold, now I know that there is no God in all the earth, but in Israel: now therefore, I pray thee, take a blessing of thy servant.

rum Dei cum uniuerso comitatu suo uenit et stetit coram eo et ait uere scio quod non sit Deus in uniuersa terra nisi tantum in Israhel obsecro itaque ut accipias benedictionem a

5:16 But he said, As the LORD liveth, before whom I seruo tuo 5:16 at ille respondit uiuit Dominus ante quem

stand, I will receive none. And he urged him to take it; but he refused.

sto quia non accipiam cumque uim faceret penitus non ad-

5:17 And Naaman said, Shall there not then, I pray thee, quieuit 5:17 dixitque Naaman ut uis sed obsecro concede

be given to thy servant two mules’ burden of earth? for thy servant will henceforth offer neither burnt offering nor

sacrifice unto other gods, but unto the LORD. mihi seruo tuo ut tollam onus duorum burdonum de terra

non enim faciet ultra seruus tuus holocaustum aut uictimam

5:18 In this thing the LORD pardon thy servant, that when diis alienis nisi Domino 5:18 hoc autem solum est de quo my master goeth into the house of Rimmon to worship there, and he leaneth on my hand, and I bow myself in the house of Rimmon: when I bow down myself in the house of Rimmon, the LORD pardon thy servant in this thing.

depreceris Dominum pro seruo tuo quando ingreditur dominus meus templum Remmon ut adoret et illo innitente super manum meam si adorauero in templo Remmon adorante me in eodem loco ut ignoscat mihi Dominus seruo tuo pro hac

5:19 And he said unto him, Go in peace. So he departed re 5:19 qui dixit ei uade in pace abiit ergo ab eo electo ter-

from him a little way.

5:20 But Gehazi, the servant of Elisha the man of God, rae tempore 5:20 dixitque Giezi puer uiri Dei pepercit do-

said, Behold, my master hath spared Naaman this Syrian, in not receiving at his hands that which he brought: but, as the LORD liveth, I will run after him, and take somewhat of him.

Liber II Regum

minus meus Naaman Syro isti ut non acciperet ab eo quae adtulit uiuit Dominus quia curram post eum et accipiam ab eo aliquid 5:21 et secutus est Giezi post tergum Naaman 5:21 So Gehazi followed after Naaman. And when Naa-

man saw him running after him, he lighted down from the chariot to meet him, and said, Is all well?

quem cum uidisset ille currentem ad se desiliuit de curru in occursum eius et ait rectene sunt omnia 5:22 et ille ait recte 5:22 And he said, All is well. My master hath sent me,

saying, Behold, even now there be come to me from mount Ephraim two young men of the sons of the prophets: give them, I pray thee, a talent of silver, and two changes of garments.

dominus meus misit me dicens modo uenerunt ad me duo adulescentes de monte Ephraim ex filiis prophetarum da eis talentum argenti et uestes mutatorias duplices 5:23 dixit- 5:23 And Naaman said, Be content, take two talents. And

he urged him, and bound two talents of silver in two bags, with two changes of garments, and laid them upon two of

que Naaman melius est ut accipias duo talenta et coegit eum his servants; and they bare them before him. ligauitque duo talenta argenti in duobus saccis et duplicia uestimenta et inposuit duobus pueris suis qui et portauerunt coram eo 5:24 cumque uenisset iam uesperi tulit de manu 5:24 And when he came to the tower, he took them from

their hand, and bestowed them in the house: and he let the men go, and they departed.

eorum et reposuit in domo dimisitque uiros et abierunt 5:25

5:25 But he went in, and stood before his master. And Elisha said unto him, Whence comest thou, Gehazi? And he said, Thy servant went no whither.

ipse autem ingressus stetit coram domino suo et dixit Heliseus unde uenis Giezi qui respondit non iuit seruus tuus quoquam 5:26 at ille nonne ait cor meum in praesenti erat 5:26 And he said unto him, Went not mine heart with thee,

when the man turned again from his chariot to meet thee? Is it a time to receive money, and to receive garments, and oliveyards, and vineyards, and sheep, and oxen, and menservants, and maidservants?

quando reuersus est homo de curru suo in occursum tui nunc igitur accepisti argentum et accepisti uestes ut emas oliueta et uineta et oues et boues et seruos et ancillas 5:27 sed et 5:27 The leprosy therefore of Naaman shall cleave unto

thee, and unto thy seed for ever. And he went out from his presence a leper as white as snow.

lepra Naaman adherebit tibi et semini tuo in sempiternum et egressus est ab eo leprosus quasi nix

6:1 dixerunt autem filii prophetarum ad Heliseum ecce lo- 6:1 And the sons of the prophets said unto Elisha, Behold

now, the place where we dwell with thee is too strait for us.

cus in quo habitamus coram te angustus est nobis 6:2 ea-

6:2 Let us go, we pray thee, unto Jordan, and take thence every man a beam, and let us make us a place there, where we may dwell. And he answered, Go ye.

mus usque ad Iordanem et tollant singuli de silua materias singulas ut aedificemus nobis ibi locum ad habitandum qui dixit ite 6:3 et ait unus ex illis ueni ergo et tu cum seruis tuis 6:3 And one said, Be content, I pray thee, and go with thy

servants. And he answered, I will go.

respondit ego ueniam 6:4 et abiit cum eis cumque uenissent 6:4 So he went with them. And when they came to Jordan,

they cut down wood.

ad Iordanem caedebant ligna 6:5 accidit autem ut cum unus 6:5 But as one was felling a beam, the axe head fell into

the water: and he cried, and said, Alas, master! for it was borrowed.

materiem succidisset caderet ferrum securis in aquam exclamauitque ille et ait eheu eheu eheu domine mi et hoc ipsum mutuo acceperam 6:6 dixit autem homo Dei ubi cecidit at 6:6 And the man of God said, Where fell it? And he

shewed him the place. And he cut down a stick, and cast it in thither; and the iron did swim.

ille monstrauit ei locum praecidit ergo lignum et misit illuc natauitque ferrum 6:7 et ait tolle qui extendit manum 6:7 Therefore said he, Take it up to thee. And he put out

his hand, and took it.

Biblia Sacra Vulgata

6:8 Then the king of Syria warred against Israel, and took et tulit illud 6:8 rex autem Syriae pugnabat contra Israhel

counsel with his servants, saying, In such and such a place shall be my camp.

consiliumque iniit cum seruis suis dicens in loco illo et illo

6:9 And the man of God sent unto the king of Israel, say- ponamus insidias 6:9 misit itaque uir Dei ad regem Israhel

ing, Beware that thou pass not such a place; for thither the Syrians are come down.

dicens caue ne transeas in loco illo quia ibi Syri in insidiis

6:10 And the king of Israel sent to the place which the man sunt 6:10 misit rex Israhel ad locum quem dixerat ei uir of God told him and warned him of, and saved himself there, not once nor twice.

Dei et praeoccupauit eum et obseruauit se ibi non semel ne-

6:11 Therefore the heart of the king of Syria was sore trou- que bis 6:11 conturbatumque est cor regis Syriae pro hac re bled for this thing; and he called his servants, and said unto them, Will ye not shew me which of us is for the king of

Israel? et conuocatis seruis suis ait quare non indicastis mihi quis 6:12 And one of his servants said, None, my lord, O king: prodiior mei sit apud regem Israhel 6:12 dixitque unus ser-

but Elisha, the prophet that is in Israel, telleth the king of Israel the words that thou speakest in thy bedchamber.

uorum eius nequaquam domine mi rex sed Heliseus propheta qui est in Israhel indicat regi Israhel omnia uerba quaecum-

6:13 And he said, Go and spy where he is, that I may send que locutus fueris in conclaui tuo 6:13 dixit eis ite et uidete

and fetch him. And it was told him, saying, Behold, he is in Dothan.

ubi sit ut mittam et capiam eum adnuntiaueruntque ei dicen-

6:14 Therefore sent he thither horses, and chariots, and a tes ecce in Dothan 6:14 misit ergo illuc equos et currus et great host: and they came by night, and compassed the city about.

robur exercitus qui cum uenissent nocte circumdederunt ci-

6:15 And when the servant of the man of God was risen uitatem 6:15 consurgens autem diluculo minister uiri Dei

early, and gone forth, behold, an host compassed the city both with horses and chariots. And his servant said unto

him, Alas, my master! how shall we do? egressus est uiditque exercitum in circuitu ciuitatis et equos

et currus nuntiauitque ei dicens eheu eheu domine mi quid

6:16 And he answered, Fear not: for they that be with us faciemus 6:16 at ille respondit noli timere plures enim no-

are more than they that be with them.

6:17 And Elisha prayed, and said, LORD, I pray thee, biscum sunt quam cum illis 6:17 cumque orasset Heliseus

open his eyes, that he may see. And the LORD opened the eyes of the young man; and he saw: and, behold, the mountain was full of horses and chariots of fire round about Elisha.

ait Domine aperi oculos huius ut uideat et aperuit Dominus oculos pueri et uidit et ecce mons plenus equorum et cur-

6:18 And when they came down to him, Elisha prayed ruum igneorum in circuitu Helisei 6:18 hostes uero descen-

unto the LORD, and said, Smite this people, I pray thee, with blindness. And he smote them with blindness accord-

ing to the word of Elisha. derunt ad eum porro Heliseus orauit Dominum dicens per-

cute obsecro gentem hanc caecitate percussitque eos Domi-

6:19 And Elisha said unto them, This is not the way, nei- nus ne uiderent iuxta uerbum Helisei 6:19 dixit autem ad

ther is this the city: follow me, and I will bring you to the man whom ye seek. But he led them to Samaria.

eos Heliseus non est haec uia nec ista est ciuitas sequimini me et ostendam uobis uirum quem quaeritis duxit ergo eos

6:20 And it came to pass, when they were come into in Samariam 6:20 cumque ingressi fuissent in Samaria dixit

Samaria, that Elisha said, LORD, open the eyes of these men, that they may see. And the LORD opened their eyes, and they saw; and, behold, they were in the midst of Samaria.

Heliseus Domine aperi oculos istorum ut uideant aperuitque Dominus oculos eorum et uiderunt esse se in medio Sama-

6:21 And the king of Israel said unto Elisha, when he saw riae 6:21 dixitque rex Israhel ad Heliseum cum uidisset eos

them, My father, shall I smite them? shall I smite them?

6:22 And he answered, Thou shalt not smite them: numquid percutiam eos pater mi 6:22 at ille ait non percu-

wouldest thou smite those whom thou hast taken captive with thy sword and with thy bow? set bread and water before them, that they may eat and drink, and go to their master.

Liber II Regum

ties neque enim cepisti eos gladio et arcu tuo ut percutias pone panem et aquam coram eis ut comedant et bibant et uadant ad dominum suum 6:23 adpositaque est eis ciborum 6:23 And he prepared great provision for them: and when

they had eaten and drunk, he sent them away, and they went to their master. So the bands of Syria came no more

magna praeparatio et comederunt et biberunt et dimisit eos into the land of Israel. abieruntque ad dominum suum et ultra non uenerunt latrones Syriae in terram Israhel 6:24 factum est autem post haec 6:24 And it came to pass after this, that Benhadad king

of Syria gathered all his host, and went up, and besieged Samaria.

congregauit Benadad rex Syriae uniuersum exercitum suum et ascendit et obsidebat Samariam 6:25 factaque est fames 6:25 And there was a great famine in Samaria: and, be-

hold, they besieged it, until an ass’s head was sold for fourscore pieces of silver, and the fourth part of a cab of

magna in Samaria et tamdiu obsessa est donec uenundare- dove’s dung for five pieces of silver. tur caput asini octoginta argenteis et quarta pars cabi stercoris columbarum quinque argenteis 6:26 cumque rex Israhel 6:26 And as the king of Israel was passing by upon the

wall, there cried a woman unto him, saying, Help, my lord, O king.

transiret per murum mulier exclamauit ad eum dicens salua me domine mi rex 6:27 qui ait non te saluet Dominus unde 6:27 And he said, If the LORD do not help thee, whence

shall I help thee? out of the barnfloor, or out of the winepress?

saluare te possum de area an de torculari dixitque ad eam rex quid tibi uis quae respondit 6:28 mulier ista dixit mihi 6:28 And the king said unto her, What aileth thee? And

she answered, This woman said unto me, Give thy son, that we may eat him to day, and we will eat my son to

da filium tuum ut comedamus eum hodie et filium meum morrow. comedemus cras 6:29 coximus ergo filium meum et come- 6:29 So we boiled my son, and did eat him: and I said

unto her on the next day, Give thy son, that we may eat him: and she hath hid her son.

dimus dixique ei die altera da filium tuum ut comedamus eum quae abscondit filium suum 6:30 quod cum audisset 6:30 And it came to pass, when the king heard the words

of the woman, that he rent his clothes; and he passed by upon the wall, and the people looked, and, behold, he had

rex scidit uestimenta sua et transiebat super murum uiditque sackcloth within upon his flesh. omnis populus cilicium quo uestitus erat ad carnem intrinsecus 6:31 et ait haec mihi faciat Deus et haec addat si steterit 6:31 Then he said, God do so and more also to me, if the

head of Elisha the son of Shaphat shall stand on him this day.

caput Helisei filii Saphat super eum hodie 6:32 Heliseus au-

6:32 But Elisha sat in his house, and the elders sat with him; and the king sent a man from before him: but ere the messenger came to him, he said to the elders, See ye how this son of a murderer hath sent to take away mine head? look, when the messenger cometh, shut the door, and hold him fast at the door: is not the sound of his master’s feet behind him?

tem sedebat in domo sua et senes sedebant cum eo praemisit itaque uirum et antequam ueniret nuntius ille dixit ad senes numquid scitis quod miserit filius homicidae hic ut praecidatur caput meum uidete ergo cum uenerit nuntius cludite ostium et non sinatis eum introire ecce enim sonitus pedum domini eius post eum est 6:33 et adhuc illo loquente eis ap- 6:33 And while he yet talked with them, behold, the mes-

senger came down unto him: and he said, Behold, this evil is of the LORD; what should I wait for the LORD any

paruit nuntius qui ueniebat ad eum et ait ecce tantum malum longer? a Domino est quid amplius expectabo a Domino

7:1 dixit autem Heliseus audite uerbum Domini haec di- 7:1 Then Elisha said, Hear ye the word of the LORD; Thus

saith the LORD, To morrow about this time shall a measure of fine flour be sold for a shekel, and two measures of

cit Dominus in tempore hoc cras modius similae uno statere barley for a shekel, in the gate of Samaria.

Biblia Sacra Vulgata

7:2 Then a lord on whose hand the king leaned answered erit et duo modii hordei statere uno in porta Samariae 7:2

the man of God, and said, Behold, if the LORD would make windows in heaven, might this thing be? And he said, Behold, thou shalt see it with thine eyes, but shalt not eat thereof.

respondens unus de ducibus super cuius manum rex incumbebat homini Dei ait si Dominus fecerit etiam cataractas in caelo numquid poterit esse quod loqueris qui ait uidebis ocu-

7:3 And there were four leprous men at the entering in of lis tuis et inde non comedes 7:3 quattuor ergo uiri erant le-

the gate: and they said one to another, Why sit we here until we die?

prosi iuxta introitum portae qui dixerunt ad inuicem quid hic

7:4 If we say, We will enter into the city, then the famine esse uolumus donec moriamur 7:4 siue ingredi uoluerimus

is in the city, and we shall die there: and if we sit still here, we die also. Now therefore come, and let us fall unto the host of the Syrians: if they save us alive, we shall live; and if they kill us, we shall but die.

ciuitatem fame moriemur siue manserimus hic moriendum nobis est uenite igitur et transfugiamus ad castra Syriae si pepercerint nobis uiuemus si autem occidere uoluerint nihilo-

7:5 And they rose up in the twilight, to go unto the camp minus moriemur 7:5 surrexerunt igitur uesperi ut uenirent

of the Syrians: and when they were come to the uttermost part of the camp of Syria, behold, there was no man there.

ad castra Syriae cumque uenissent ad principium castrorum

7:6 For the LORD had made the host of the Syrians to hear Syriae nullum ibidem reppererunt 7:6 siquidem Dominus

a noise of chariots, and a noise of horses, even the noise of a great host: and they said one to another, Lo, the king of Israel hath hired against us the kings of the Hittites, and the kings of the Egyptians, to come upon us.

sonitum audiri fecerat in castris Syriae curruum et equorum et exercitus plurimi dixeruntque ad inuicem ecce mercede conduxit aduersum nos rex Israhel reges Hettheorum et Ae-

7:7 Wherefore they arose and fled in the twilight, and left gyptiorum et uenerunt super nos 7:7 surrexerunt ergo et fu-

their tents, and their horses, and their asses, even the camp as it was, and fled for their life.

gerunt in tenebris et dereliquerunt tentoria sua et equos et asinos in castris fugeruntque animas tantum suas saluare cu-

7:8 And when these lepers came to the uttermost part of pientes 7:8 igitur cum uenissent leprosi illi ad principium

the camp, they went into one tent, and did eat and drink, and carried thence silver, and gold, and raiment, and went and hid it; and came again, and entered into another tent, and carried thence also, and went and hid it.

castrorum ingressi sunt unum tabernaculum et comederunt et biberunt tuleruntque inde argentum et aurum et uestes et abierunt et absconderunt et rursum reuersi sunt ad aliud ta-

7:9 Then they said one to another, We do not well: this bernaculum et inde similiter auferentes absconderunt 7:9

day is a day of good tidings, and we hold our peace: if we tarry till the morning light, some mischief will come upon us: now therefore come, that we may go and tell the king’s household.

dixeruntque ad inuicem non recte facimus haec enim dies boni nuntii est si tacuerimus et noluerimus nuntiare usque mane sceleris arguemur uenite eamus et nuntiemus in aula

7:10 So they came and called unto the porter of the city: regis 7:10 cumque uenissent ad portam ciuitatis narraue-

and they told them, saying, We came to the camp of the Syrians, and, behold, there was no man there, neither voice of man, but horses tied, and asses tied, and the tents as they were.

runt eis dicentes iuimus ad castra Syriae et nullum ibidem repperimus hominum nisi equos et asinos alligatos et fixa

7:11 And he called the porters; and they told it to the king’s tentoria 7:11 ierunt ergo portarii et nuntiauerunt in palatio

house within.

7:12 And the king arose in the night, and said unto his ser- regis intrinsecus 7:12 qui surrexit nocte et ait ad seruos suos

vants, I will now shew you what the Syrians have done to us. They know that we be hungry; therefore are they gone out of the camp to hide themselves in the field, saying, When they come out of the city, we shall catch them alive, and get into the city.

dico uobis quid fecerint nobis Syri sciunt quia fame labora-

Liber II Regum

mus et idcirco egressi sunt de castris et latitant in agris dicentes cum egressi fuerint de ciuitate capiemus eos uiuentes et tunc ciuitatem ingredi poterimus 7:13 respondit autem 7:13 And one of his servants answered and said, Let some

take, I pray thee, five of the horses that remain, which are left in the city, (behold, they are as all the multitude of Israel that are left in it: behold, I say, they are even as all the multitude of the Israelites that are consumed:) and let us send and see.

unus seruorum eius tollamus quinque equos qui remanserunt in urbe quia ipsi tantum sunt in uniuersa multitudine Israhel alii enim consumpti sunt et mittentes explorare poterimus 7:14 adduxerunt ergo duos equos misitque rex ad castra Sy- 7:14 They took therefore two chariot horses; and the king

sent after the host of the Syrians, saying, Go and see.

rorum dicens ite uidete 7:15 qui abierunt post eos usque ad 7:15 And they went after them unto Jordan: and, lo, all the

way was full of garments and vessels, which the Syrians had cast away in their haste. And the messengers returned,

Iordanem ecce autem omnis uia plena erat uestibus et ua- and told the king. sis quae proiecerant Syri cum turbarentur reuersique nuntii indicauerunt regi 7:16 et egressus populus diripuit castra 7:16 And the people went out, and spoiled the tents of the

Syrians. So a measure of fine flour was sold for a shekel, and two measures of barley for a shekel, according to the

Syriae factusque est modius similae statere uno et duo mo- word of the LORD. dii hordei statere uno iuxta uerbum Domini 7:17 porro rex 7:17 And the king appointed the lord on whose hand he

leaned to have the charge of the gate: and the people trode upon him in the gate, and he died, as the man of God had

ducem illum in cuius manu incubuerat constituit ad portam said, who spake when the king came down to him. quem conculcauit turba in introitu et mortuus est iuxta quod locutus fuerat uir Dei quando descenderat rex ad eum 7:18 7:18 And it came to pass as the man of God had spoken

to the king, saying, Two measures of barley for a shekel, and a measure of fine flour for a shekel, shall be to morrow

factumque est secundum sermonem uiri Dei quem dixerat about this time in the gate of Samaria: regi quando ait duo modii hordei statere uno erunt et modius similae statere uno hoc eodem tempore cras in porta Samariae 7:19 quando responderat dux ille uiro Dei et dixerat 7:19 And that lord answered the man of God, and said,

Now, behold, if the LORD should make windows in heaven, might such a thing be? And he said, Behold, thou

etiam si Dominus fecerit cataractas in caelo numquid fieri shalt see it with thine eyes, but shalt not eat thereof. poterit quod loqueris et dixit ei uidebis oculis tuis et inde non comedes 7:20 euenit ergo ei sicut praedictum erat et 7:20 And so it fell out unto him: for the people trode upon

him in the gate, and he died.

conculcauit eum populus in porta et mortuus est

8:1 Heliseus autem locutus est ad mulierem cuius uiuere 8:1 Then spake Elisha unto the woman, whose son he

had restored to life, saying, Arise, and go thou and thine household, and sojourn wheresoever thou canst sojourn: for the LORD hath called for a famine; and it shall also come upon the land seven years.

fecerat filium dicens surge uade tu et domus tua et peregrinare ubicumque reppereris uocauit enim Dominus famem et ueniet super terram septem annis 8:2 quae surrexit et fe- 8:2 And the woman arose, and did after the saying of the

man of God: and she went with her household, and sojourned in the land of the Philistines seven years.

cit iuxta uerbum hominis Dei et uadens cum domo sua peregrinata est in terra Philisthim diebus multis 8:3 cumque 8:3 And it came to pass at the seven years’ end, that the

woman returned out of the land of the Philistines: and she went forth to cry unto the king for her house and for her

finiti essent anni septem reuersa est mulier de terra Philis- land. thim et egressa est ut interpellaret regem pro domo sua et agris suis 8:4 rex autem loquebatur cum Giezi puero uiri 8:4 And the king talked with Gehazi the servant of the man

of God, saying, Tell me, I pray thee, all the great things that Elisha hath done.

Biblia Sacra Vulgata

Dei dicens narra mihi omnia magnalia quae fecit Heliseus

8:5 And it came to pass, as he was telling the king how he 8:5 cumque ille narraret regi quomodo mortuum suscitasset

had restored a dead body to life, that, behold, the woman, whose son he had restored to life, cried to the king for her house and for her land. And Gehazi said, My lord, O king, this is the woman, and this is her son, whom Elisha restored to life.

apparuit mulier cuius uiuificauerat filium clamans ad regem pro domo sua et pro agris suis dixitque Giezi domine mi rex haec est mulier et hic filius eius quem suscitauit Heliseus

8:6 And when the king asked the woman, she told him. 8:6 et interrogauit rex mulierem quae narrauit ei deditque ei

So the king appointed unto her a certain officer, saying, Restore all that was hers, and all the fruits of the field since

the day that she left the land, even until now. rex eunuchum unum dicens restitue ei omnia quae sua sunt et

uniuersos rediius agrorum a die qua reliquit terram usque ad

8:7 And Elisha came to Damascus; and Benhadad the king praesens 8:7 uenit quoque Heliseus Damascum et Benadad

of Syria was sick; and it was told him, saying, The man of God is come hither.

rex Syriae aegrotabat nuntiaueruntque ei dicentes uenit uir

8:8 And the king said unto Hazael, Take a present in thine Dei huc 8:8 et ait rex ad Azahel tolle tecum munera et uade hand, and go, meet the man of God, and enquire of the LORD by him, saying, Shall I recover of this disease?

in occursum uiri Dei et consule Dominum per eum dicens si

8:9 So Hazael went to meet him, and took a present with euadere potero de infirmitate mea hac 8:9 iuit igitur Azahel

him, even of every good thing of Damascus, forty camels’ burden, and came and stood before him, and said, Thy son Benhadad king of Syria hath sent me to thee, saying, Shall I recover of this disease?

in occursum eius habens secum munera et omnia bona Damasci onera quadraginta camelorum cumque stetisset coram eo ait filius tuus Benadad rex Syriae misit me ad te dicens si

8:10 And Elisha said unto him, Go, say unto him, Thou sanari potero de infirmitate mea hac 8:10 dixitque ei Heli-

mayest certainly recover: howbeit the LORD hath shewed me that he shall surely die.

seus uade dic ei sanaberis porro ostendit mihi Dominus quia

8:11 And he settled his countenance stedfastly, until he morte morietur 8:11 stetitque cum eo et conturbatus est us-

was ashamed: and the man of God wept.

8:12 And Hazael said, Why weepeth my lord? And he que ad suffusionem uultus fleuitque uir Dei 8:12 cui Aza-

answered, Because I know the evil that thou wilt do unto the children of Israel: their strong holds wilt thou set on fire, and their young men wilt thou slay with the sword, and wilt dash their children, and rip up their women with child.

hel ait quare dominus meus flet at ille respondit quia scio quae facturus sis filiis Israhel mala ciuitates eorum munitas igne succendes et iuuenes eorum interficies gladio et paruu-

8:13 And Hazael said, But what, is thy servant a dog, that los eorum elides et praegnantes diuides 8:13 dixitque Aza-

he should do this great thing? And Elisha answered, The LORD hath shewed me that thou shalt be king over Syria.

hel quid enim sum seruus tuus canis ut faciam rem istam magnam et ait Heliseus ostendit mihi Dominus te regem Syriae

8:14 So he departed from Elisha, and came to his mas- fore 8:14 qui cum recessisset ab Heliseo uenit ad dominum

ter; who said to him, What said Elisha to thee? And he answered, He told me that thou shouldest surely recover.

suum qui ait ei quid tibi dixit Heliseus at ille respondit dixit

8:15 And it came to pass on the morrow, that he took a mihi recipiet sanitatem 8:15 cumque uenisset dies altera tu-

thick cloth, and dipped it in water, and spread it on his face, so that he died: and Hazael reigned in his stead.

lit sagulum et infudit aqua et expandit super faciem eius quo

8:16 And in the fifth year of Joram the son of Ahab king mortuo regnauit Azahel pro eo 8:16 anno quinto Ioram fi-

of Israel, Jehoshaphat being then king of Judah, Jehoram the son of Je hoshaphat king of Judah began to reign.

lii Ahab regis Israhel et Iosaphat regis Iuda regnauit Ioram

8:17 Thirty and two years old was he when he began to filius Iosaphat rex Iuda 8:17 triginta duorum erat annorum

reign; and he reigned eight years in Jerusalem.

Liber II Regum

cum regnare coepisset et octo annis regnauit in Hierusalem 8:18 ambulauitque in uiis regum Israhel sicut ambulauerat 8:18 And he walked in the way of the kings of Israel, as

did the house of Ahab: for the daughter of Ahab was his wife: and he did evil in the sight of the LORD.

domus Ahab filia enim Ahab erat uxor eius et fecit quod malum est coram Domino 8:19 noluit autem Dominus dis- 8:19 Yet the LORD would not destroy Judah for David his

servant’s sake, as he promised him to give him alway a light, and to his children.

perdere Iudam propter Dauid seruum suum sicut promiserat ei ut daret illi lucernam et filiis eius cunctis diebus 8:20 in 8:20 In his days Edom revolted from under the hand of

Judah, and made a king over themselves.

diebus eius recessit Edom ne esset sub Iuda et constituit sibi

regem 8:21 uenitque Ioram Seira et omnis currus cum eo 8:21 So Joram went over to Zair, and all the chariots with him: and he rose by night, and smote the Edomites which compassed him about, and the captains of the chariots: and

et surrexit nocte percussitque Idumeos qui eum circumde- the people fled into their tents. derant et principes curruum populus autem fugit in tabernacula sua 8:22 recessit ergo Edom ne esset sub Iuda usque 8:22 Yet Edom revolted from under the hand of Judah unto

this day. Then Libnah revolted at the same time.

ad diem hanc tunc recessit et Lobna in tempore illo 8:23 8:23 And the rest of the acts of Joram, and all that he did, are they not written in the book of the chronicles of the kings of Judah?

reliqua autem sermonum Ioram et uniuersa quae fecit nonne haec scripta sunt in libro uerborum dierum regum Iuda 8:24 8:24 And Joram slept with his fathers, and was buried

with his fathers in the city of David: and Ahaziah his son reigned in his stead.

et dormiuit Ioram cum patribus suis sepultusque est cum eis in ciuitate Dauid et regnauit Ahazias filius eius pro eo 8:25 8:25 In the twelfth year of Joram the son of Ahab king of

Israel did Ahaziah the son of Jehoram king of Judah begin to reign.

anno duodecimo Ioram filii Ahab regis Israhel regnauit Ahazias filius Ioram regis Iudae 8:26 uiginti duorum annorum 8:26 Two and twenty years old was Ahaziah when he be-

gan to reign; and he reigned one year in Jerusalem. And his mother’s name was Athaliah, the daughter of Omri

erat Ahazias cum regnare coepisset et uno anno regnauit in king of Israel. Hierusalem nomen matris eius Athalia filia Amri regis Israhel 8:27 et ambulauit in uiis domus Ahab et fecit quod ma- 8:27 And he walked in the way of the house of Ahab,

and did evil in the sight of the LORD, as did the house of Ahab: for he was the son in law of the house of Ahab.

lum est coram Domino sicut domus Ahab gener enim domus Ahab fuit 8:28 abiit quoque cum Ioram filio Ahab ad pro- 8:28 And he went with Joram the son of Ahab to the war

against Hazael king of Syria in Ramothgilead; and the Syrians wounded Joram.

eliandum contra Azahel regem Syriae in Ramoth Galaad et uulnerauerunt Syri Ioram 8:29 qui reuersus est ut curaretur 8:29 And king Joram went back to be healed in Jezreel of

the wounds which the Syrians had given him at Ramah, when he fought against Hazael king of Syria. And Ahaziah the son of Jehoram king of Judah went down to see Joram the son of Ahab in Jezreel, because he was sick.

in Hiezrahel quia uulnerauerant eum Syri in Rama proeliantem contra Azahel regem Syriae porro Ahazias filius Ioram rex Iuda descendit inuisere Ioram filium Ahab in Hiezrahel quia aegrotabat

9:1 Heliseus autem prophetes uocauit unum de filiis pro- 9:1 And Elisha the prophet called one of the children of

the prophets, and said unto him, Gird up thy loins, and take this box of oil in thine hand, and go to Ramothgilead:

phetarum et ait illi accinge lumbos tuos et tolle lenticulam

olei hanc in manu tua et uade in Ramoth Galaad 9:2 cum- 9:2 And when thou comest thither, look out there Jehu the son of Jehoshaphat the son of Nimshi, and go in, and make him arise up from among his brethren, and carry him to an

que ueneris illuc uidebis Hieu filium Iosaphat filii Namsi inner chamber;

Biblia Sacra Vulgata

et ingressus suscitabis eum de medio fratrum suorum et in-

9:3 Then take the box of oil, and pour it on his head, and troduces interius cubiculum 9:3 tenensque lenticulam olei

say, Thus saith the LORD, I have anointed thee king over Israel. Then open the door, and flee, and tarry not.

fundes super caput eius et dices haec dicit Dominus unxi te regem super Israhel aperiesque ostium et fugies et non ibi

9:4 So the young man, even the young man the prophet, subsistes 9:4 abiit ergo adulescens puer prophetae Ramoth

went to Ramothgilead.

9:5 And when he came, behold, the captains of the host Galaad 9:5 et ingressus est ecce autem principes exercitus

were sitting; and he said, I have an errand to thee, O captain. And Jehu said, Unto which of all us? And he said,

To thee, O captain. sedebant et ait uerbum mihi ad te princeps dixitque Hieu ad

9:6 And he arose, and went into the house; and he poured quem ex omnibus nobis at ille dixit ad te o princeps 9:6 the oil on his head, and said unto him, Thus saith the LORD God of Israel, I have anointed thee king over the

people of the LORD, even over Israel. et surrexit et ingressus est cubiculum at ille fudit oleum su-

per caput eius et ait haec dicit Dominus Deus Israhel unxi te

9:7 And thou shalt smite the house of Ahab thy master, regem super populum Domini Israhel 9:7 et percuties do-

that I may avenge the blood of my servants the prophets, and the blood of all the servants of the LORD, at the hand

of Jezebel. mum Ahab domini tui ut ulciscar sanguinem seruorum me-

orum prophetarum et sanguinem omnium seruorum Domini

9:8 For the whole house of Ahab shall perish: and I will de manu Hiezabel 9:8 perdamque omnem domum Ahab et cut off from Ahab him that pisseth against the wall, and him that is shut up and left in Israel:

interficiam de Ahab mingentem ad parietem et clausum et

9:9 And I will make the house of Ahab like the house of nouissimum in Israhel 9:9 et dabo domum Ahab sicut do-

Jeroboam the son of Nebat, and like the house of Baasha the son of Ahijah:

mum Hieroboam filii Nabath et sicut domum Baasa filii Ahia

9:10 And the dogs shall eat Jezebel in the portion of 9:10 Hiezabel quoque comedent canes in agro Hiezrahel nec

Jezreel, and there shall be none to bury her. And he opened the door, and fled.

erit qui sepeliat eam aperuitque ostium et fugit 9:11 Hieu

9:11 Then Jehu came forth to the servants of his lord: and one said unto him, Is all well? wherefore came this mad fellow to thee? And he said unto them, Ye know the man, and his communication.

autem egressus est ad seruos domini sui qui dixerunt ei rectene sunt omnia quid uenit insanus iste ad te qui ait eis no-

9:12 And they said, It is false; tell us now. And he said, stis hominem et quid locutus sit 9:12 at illi responderunt

Thus and thus spake he to me, saying, Thus saith the LORD, I have anointed thee king over Israel.

falsum est sed magis narra nobis qui ait eis haec et haec locutus est mihi et ait haec dicit Dominus unxi te regem su-

9:13 Then they hasted, and took every man his garment, per Israhel 9:13 festinauerunt itaque et unusquisque tollens

and put it under him on the top of the stairs, and blew with trumpets, saying, Jehu is king.

pallium suum posuerunt sub pedibus eius in similitudinem tribunalis et cecinerunt tuba atque dixerunt regnauit Hieu

9:14 So Jehu the son of Jehoshaphat the son of Nimshi 9:14 coniurauit ergo Hieu filius Iosaphat filii Namsi contra

conspired against Joram. (Now Joram had kept Ramothgilead, he and all Israel, because of Hazael king of Syria.

Ioram porro Ioram obsederat Ramoth Galaad ipse et omnis

9:15 But king Joram was returned to be healed in Jezreel Israhel contra Azahel regem Syriae 9:15 et reuersus fuerat

of the wounds which the Syrians had given him, when he fought with Hazael king of Syria.) And Jehu said, If it be your minds, then let none go forth nor escape out of the city to go to tell it in Jezreel.

ut curaretur in Hiezrahel propter uulnera quia percusserant eum Syri proeliantem contra Azahel regem Syriae dixitque Hieu si placet uobis nemo egrediatur profugus de ciuitate ne

Liber II Regum

uadat et nuntiet in Hiezrahel 9:16 et ascendit et profectus 9:16 So Jehu rode in a chariot, and went to Jezreel; for

Joram lay there. And Ahaziah king of Judah was come down to see Joram.

est in Hiezrahel Ioram enim aegrotabat ibi et Ahazia rex Iuda descenderat ad uisitandum Ioram 9:17 igitur speculator qui 9:17 And there stood a watchman on the tower in Jezreel,

and he spied the company of Jehu as he came, and said, I see a company. And Joram said, Take an horseman, and

stabat super turrem Hiezrahel uidit globum Hieu uenientis et send to meet them, and let him say, Is it peace? ait uideo ego globum dixitque Ioram tolle currum et mitte in occursum eorum et dicat uadens rectene sunt omnia 9:18 9:18 So there went one on horseback to meet him, and

said, Thus saith the king, Is it peace? And Jehu said, What hast thou to do with peace? turn thee behind me. And the watchman told, saying, The messenger came to them, but he cometh not again.

abiit igitur qui ascenderat currum in occursum eius et ait haec dicit rex pacata sunt omnia dixitque ei Hieu quid tibi et paci transi et sequere me nuntiauit quoque speculator dicens uenit nuntius ad eos et non reuertitur 9:19 misit etiam 9:19 Then he sent out a second on horseback, which came

to them, and said, Thus saith the king, Is it peace? And Jehu answered, What hast thou to do with peace? turn

currum equorum secundum uenitque ad eos et ait haec dicit thee behind me. rex num pax est et ait Hieu quid tibi et paci transi et sequere me 9:20 nuntiauit autem speculator dicens uenit usque ad 9:20 And the watchman told, saying, He came even unto

them, and cometh not again: and the driving is like the driving of Jehu the son of Nimshi; for he driveth furiously.

eos et non reuertitur est autem incessus quasi incessus Hieu filii Namsi praeceps enim graditur 9:21 et ait Ioram iunge 9:21 And Joram said, Make ready. And his chariot was

made ready. And Joram king of Israel and Ahaziah king of Judah went out, each in his chariot, and they went out against Jehu, and met him in the portion of Naboth the Jezreelite.

currum iunxeruntque currum eius et egressus est Ioram rex Israhel et Ahazias rex Iuda singuli in curribus suis egressique sunt in occursum Hieu et inuenerunt eum in agro Naboth Hiezrahelitis 9:22 cumque uidisset Ioram Hieu dixit pax est 9:22 And it came to pass, when Joram saw Jehu, that he

said, Is it peace, Jehu? And he answered, What peace, so long as the whoredoms of thy mother Jezebel and her

Hieu at ille respondit quae pax adhuc fornicationes Hiezabel witchcrafts are so many? matris tuae et ueneficia eius multa uigent 9:23 conuertit au- 9:23 And Joram turned his hands, and fled, and said to

Ahaziah, There is treachery, O Ahaziah.

tem Ioram manum suam et fugiens ait ad Ahaziam insidiae Ahazia 9:24 porro Hieu tetendit arcum manu et percussit 9:24 And Jehu drew a bow with his full strength, and

smote Jehoram between his arms, and the arrow went out at his heart, and he sunk down in his chariot.

Ioram inter scapulas et egressa est sagitta per cor eius statimque corruit in curru suo 9:25 dixitque Hieu ad Baddacer 9:25 Then said Jehu to Bidkar his captain, Take up, and

cast him in the portion of the field of Naboth the Jezreelite: for remember how that, when I and thou rode together

ducem tolle proice eum in agro Naboth Hiezrahelitae me- after Ahab his father, the LORD laid this burden upon him;

mini enim quando ego et tu sedentes in curru sequebamur Ahab patrem huius quod Dominus onus hoc leuauerit super

eum dicens 9:26 si non pro sanguine Naboth et pro san- 9:26 Surely I have seen yesterday the blood of Naboth, and the blood of his sons, saith the LORD; and I will requite thee in this plat, saith the LORD. Now therefore take and cast him into the plat of ground, according to the word of the LORD.

guine filiorum eius quem uidi heri ait Dominus reddam tibi in agro isto dicit Dominus nunc igitur tolle proice eum in agro iuxta uerbum Domini 9:27 Ahazias autem rex Iuda ui- 9:27 But when Ahaziah the king of Judah saw this, he fled

by the way of the garden house. And Jehu followed after him, and said, Smite him also in the chariot. And they did so at the going up to Gur, which is by Ibleam. And he fled to Megiddo, and died there.

dens hoc fugit per uiam domus horti persecutusque est eum

Biblia Sacra Vulgata

Hieu et ait etiam hunc percutite in curru suo in ascensu Gaber qui est iuxta Ieblaam qui fugit in Mageddo et mortuus est

9:28 And his servants carried him in a chariot to ibi 9:28 et inposuerunt eum serui eius super currum suum

Jerusalem, and buried him in his sepulchre with his fathers in the city of David.

et tulerunt Hierusalem sepelieruntque in sepulchro cum pat-

9:29 And in the eleventh year of Joram the son of Ahab ribus suis in ciuitate Dauid 9:29 anno undecimo Ioram filii

began Ahaziah to reign over Judah. 9:30 And when Jehu was come to Jezreel, Jezebel heard of Ahab rege Ahazia super Iudam 9:30 uenit Hieu Hiezrahel it; and she painted her face, and tired her head, and looked out at a window.

porro Hiezabel introitu eius audito depinxit oculos suos sti-

9:31 And as Jehu entered in at the gate, she said, Had bio et ornauit caput suum et respexit per fenestram 9:31 in-

Zimri peace, who slew his master?

gredientem Hieu per portam et ait numquid pax esse potest

9:32 And he lifted up his face to the window, and said, Zamri qui interfecit dominum suum 9:32 leuauitque Hieu

Who is on my side? who? And there looked out to him two or three eunuchs.

faciem suam ad fenestram et ait quae est ista et inclinauerunt

9:33 And he said, Throw her down. So they threw her se ad eum duo uel tres eunuchi 9:33 at ille dixit eis prae-

down: and some of her blood was sprinkled on the wall, and on the horses: and he trode her under foot.

cipitate eam deorsum et praecipitauerunt eam aspersusque est sanguine paries et equorum ungulae qui conculcauerunt

9:34 And when he was come in, he did eat and drink, and eam 9:34 cumque ingressus esset et comederet bibissetque

said, Go, see now this cursed woman, and bury her: for she is a king’s daughter.

ait ite uidete maledictam illam et sepelite eam quia filia re-

9:35 And they went to bury her: but they found no more gis est 9:35 cumque issent ut sepelirent eam non inuenerunt

of her than the skull, and the feet, and the palms of her hands.

nisi caluariam et pedes et summas manus 9:36 reuersique

9:36 Wherefore they came again, and told him. And he said, This is the word of the LORD, which he spake by his servant Elijah the Tishbite, saying, In the portion of Jezreel shall dogs eat the flesh of Jezebel:

nuntiauerunt ei et ait Hieu sermo Domini est quem locutus est per seruum suum Heliam Thesbiten dicens in agro Hiez-

9:37 And the carcase of Jezebel shall be as dung upon the rahel comedent canes carnes Hiezabel 9:37 et erunt carnes

face of the field in the portion of Jezreel; so that they shall not say, This is Jezebel.

Hiezabel sicut stercus super faciem terrae in agro Hiezrahel ita ut praetereuntes dicant haecine est illa Hiezabel

10:1 And Ahab had seventy sons in Samaria. And Jehu 10:1 erant autem Ahab septuaginta filii in Samaria scripsit

wrote letters, and sent to Samaria, unto the rulers of Jezreel, to the elders, and to them that brought up Ahab’s

children, saying, ergo Hieu litteras et misit in Samariam ad optimates ciuitatis 10:2 Now as soon as this letter cometh to you, seeing your et ad maiores natu et ad nutricios Ahab dicens 10:2 statim

master’s sons are with you, and there are with you chariots and horses, a fenced city also, and armour;

ut acceperitis litteras has qui habetis filios domini uestri et

10:3 Look even out the best and meetest of your master’s currus et equos et ciuitates firmas et arma 10:3 eligite me-

sons, and set him on his father’s throne, and fight for your master’s house.

liorem et eum qui uobis placuerit de filiis domini uestri et ponite eum super solium patris sui et pugnate pro domo do-

10:4 But they were exceedingly afraid, and said, Behold, mini uestri 10:4 timuerunt illi uehementer et dixerunt ecce

two kings stood not before him: how then shall we stand?

duo reges non potuerunt stare coram eo et quomodo nos ua-

10:5 And he that was over the house, and he that was over lebimus resistere 10:5 miserunt ergo praepositus domus et

the city, the elders also, and the bringers up of the children, sent to Jehu, saying, We are thy servants, and will do all that thou shalt bid us; we will not make any king: do thou that which is good in thine eyes.

Liber II Regum

praefectus ciuitatis et maiores natu et nutricii ad Hieu dicentes serui tui sumus quaecumque iusseris faciemus nec constituemus regem quodcumque tibi placet fac 10:6 rescrip- 10:6 Then he wrote a letter the second time to them, say-

ing, If ye be mine, and if ye will hearken unto my voice, take ye the heads of the men your master’s sons, and come to me to Jezreel by to morrow this time. Now the king’s sons, being seventy persons, were with the great men of the city, which brought them up.

sit autem eis litteras secundo dicens si mei estis et oboeditis mihi tollite capita filiorum domini uestri et uenite ad me hac eadem hora cras in Hiezrahel porro filii regis septuaginta uiri apud optimates ciuitatis nutriebantur 10:7 cumque ue- 10:7 And it came to pass, when the letter came to them,

that they took the king’s sons, and slew seventy persons, and put their heads in baskets, and sent him them to

nissent litterae ad eos tulerunt filios regis et occiderunt sep- Jezreel. tuaginta uiros et posuerunt capita eorum in cofinis et miserunt ad eum in Hiezrahel 10:8 uenit autem nuntius et indi- 10:8 And there came a messenger, and told him, saying,

They have brought the heads of the king’s sons. And he said, Lay ye them in two heaps at the entering in of the

cauit ei dicens adtulerunt capita filiorum regis qui respondit gate until the morning. ponite ea duos aceruos iuxta introitum portae usque mane 10:9 cumque diluxisset egressus est et stans dixit ad omnem 10:9 And it came to pass in the morning, that he went

out, and stood, and said to all the people, Ye be righteous: behold, I conspired against my master, and slew him: but

populum iusti estis si ego coniuraui contra dominum meum who slew all these? et interfeci eum quis percussit omnes hos 10:10 uidete ergo 10:10 Know now that there shall fall unto the earth noth-

ing of the word of the LORD, which the LORD spake concerning the house of Ahab: for the LORD hath done that

nunc quoniam non cecidit de sermonibus Domini in terram which he spake by his servant Elijah. quos locutus est Dominus super domum Ahab et Dominus fecit quod locutus est in manu serui sui Heliae 10:11 per- 10:11 So Jehu slew all that remained of the house of Ahab

in Jezreel, and all his great men, and his kinsfolks, and his priests, until he left him none remaining.

cussit igitur Hieu omnes qui reliqui erant de domo Ahab in Hiezrahel et uniuersos optimates eius et notos et sacerdotes donec non remanerent ex eo reliquiae 10:12 et surrexit 10:12 And he arose and departed, and came to Samaria.

And as he was at the shearing house in the way,

et uenit in Samariam cumque uenisset ad Camaram pastorum in uia 10:13 inuenit fratres Ahaziae regis Iuda dixitque 10:13 Jehu met with the brethren of Ahaziah king of Ju-

dah, and said, Who are ye? And they answered, We are the brethren of Ahaziah; and we go down to salute the

ad eos quinam estis uos at illi responderunt fratres Ahaziae children of the king and the children of the queen. sumus et descendimus ad salutandos filios regis et filios reginae 10:14 qui ait conprehendite eos uiuos quos cum con- 10:14 And he said, Take them alive. And they took them

alive, and slew them at the pit of the shearing house, even two and forty men; neither left he any of them.

prehendissent uiuos iugulauerunt eos in cisterna iuxta Camaram quadraginta duos uiros et non reliquit ex eis quemquam 10:15 cumque abisset inde inuenit Ionadab filium Rechab in 10:15 And when he was departed thence, he lighted on

Jehonadab the son of Rechab coming to meet him: and he saluted him, and said to him, Is thine heart right, as my heart is with thy heart? And Jehonadab answered, It is. If it be, give me thine hand. And he gave him his hand; and he took him up to him into the chariot.

occursum sibi et benedixit ei et ait ad eum numquid est cor tuum rectum sicut cor meum cum corde tuo et ait Ionadab est si est inquit da manum tuam qui dedit manum suam at ille leuauit eum ad se in curru 10:16 dixitque ad eum ueni 10:16 And he said, Come with me, and see my zeal for the

LORD. So they made him ride in his chariot.

Biblia Sacra Vulgata

mecum et uide zelum meum pro Domino et inpositum cur-

10:17 And when he came to Samaria, he slew all that re- rui suo 10:17 duxit in Samariam et percussit omnes qui re-

mained unto Ahab in Samaria, till he had destroyed him, according to the saying of the LORD, which he spake to

Elijah. liqui fuerant de Ahab in Samaria usque ad unum iuxta uer-

10:18 And Jehu gathered all the people together, and said bum Domini quod locutus est per Heliam 10:18 congregaunto them, Ahab served Baal a little; but Jehu shall serve him much.

uit ergo Hieu omnem populum et dixit ad eos Ahab coluit

10:19 Now therefore call unto me all the prophets of Baal, Baal parum ego autem colam eum amplius 10:19 nunc igiall his servants, and all his priests; let none be wanting: for I have a great sacrifice to do to Baal; whosoever shall be wanting, he shall not live. But Jehu did it in subtilty, to the intent that he might destroy the worshippers of Baal.

tur omnes prophetas Baal et uniuersos seruos eius et cunctos sacerdotes ipsius uocate ad me nullus sit qui non ueniat sacrificium enim grande est mihi Baal quicumque defuerit non uiuet porro Hieu faciebat hoc insidiose ut disperderet

10:20 And Jehu said, Proclaim a solemn assembly for cultores Baal 10:20 dixit sanctificate diem sollemnem Baal

Baal. And they proclaimed it.

10:21 And Jehu sent through all Israel: and all the wor- uocauitque 10:21 et misit in uniuersos terminos Israhel et

shippers of Baal came, so that there was not a man left that came not. And they came into the house of Baal; and

the house of Baal was full from one end to another. uenerunt cuncti serui Baal non fuit residuus ne unus qui-

dem qui non ueniret et ingressi sunt templum Baal et repleta

10:22 And he said unto him that was over the vestry, Bring est domus Baal a summo usque ad summum 10:22 dixitque forth vestments for all the worshippers of Baal. And he brought them forth vestments.

his qui erant super uestes proferte uestimenta uniuersis ser-

10:23 And Jehu went, and Jehonadab the son of Rechab, uis Baal et protulerunt eis uestes 10:23 ingressusque Hieu

into the house of Baal, and said unto the worshippers of Baal, Search, and look that there be here with you none of the servants of the LORD, but the worshippers of Baal only.

et Ionadab filius Rechab templum Baal et ait cultoribus Baal perquirite et uidete ne quis forte uobiscum sit de seruis Do-

10:24 And when they went in to offer sacrifices and burnt mini sed ut sint soli serui Baal 10:24 ingressi sunt igitur

offerings, Jehu appointed fourscore men without, and said, If any of the men whom I have brought into your hands escape, he that letteth him go, his life shall be for the life of him.

ut facerent uictimas et holocausta Hieu autem praeparauerat sibi foris octoginta uiros et dixerat eis quicumque fugerit de hominibus his quos ego adduxero in manus uestras anima

10:25 And it came to pass, as soon as he had made an end eius erit pro anima illius 10:25 factum est ergo cum con-

of offering the burnt offering, that Jehu said to the guard and to the captains, Go in, and slay them; let none come forth. And they smote them with the edge of the sword; and the guard and the captains cast them out, and went to the city of the house of Baal.

pletum esset holocaustum praecepit Hieu militibus et ducibus suis ingredimini et percutite eos nullus euadat percusseruntque eos ore gladii et proiecerunt milites et duces et

10:26 And they brought forth the images out of the house ierunt in ciuitatem templi Baal 10:26 et protulerunt statuam

of Baal, and burned them.

10:27 And they brake down the image of Baal, and brake de fano Baal et conbuserunt 10:27 et comminuerunt eam

down the house of Baal, and made it a draught house unto this day.

destruxerunt quoque aedem Baal et fecerunt pro ea latrinas

10:28 Thus Jehu destroyed Baal out of Israel. usque ad diem hanc 10:28 deleuit itaque Hieu Baal de Is-10:29 Howbeit from the sins of Jeroboam the son of Nebat, rahel 10:29 uerumtamen a peccatis Hieroboam filii Nabath

who made Israel to sin, Jehu departed not from after them, to wit, the golden calves that were in Bethel, and that were

in Dan. qui peccare fecerat Israhel non recessit nec dereliquit uitu-

Liber II Regum

los aureos qui erant in Bethel et in Dan 10:30 dixit autem 10:30 And the LORD said unto Jehu, Because thou hast

done well in executing that which is right in mine eyes, and hast done unto the house of Ahab according to all that was in mine heart, thy children of the fourth generation shall sit on the throne of Israel.

Dominus ad Hieu quia studiose fecisti quod rectum erat et placebat in oculis meis et omnia quae erant in corde meo fecisti contra domum Ahab filii tui usque ad quartam generationem sedebunt super thronum Israhel 10:31 porro Hieu 10:31 But Jehu took no heed to walk in the law of the

LORD God of Israel with all his heart: for he departed not from the sins of Jeroboam, which made Israel to sin.

non custodiuit ut ambularet in lege Domini Dei Israhel in toto corde suo non enim recessit a peccatis Hieroboam qui peccare fecerat Israhel 10:32 in diebus illis coepit Domi- 10:32 In those days the LORD began to cut Israel short:

and Hazael smote them in all the coasts of Israel;

nus taedere super Israhel percussitque eos Azahel in uniuersis finibus Israhel 10:33 a Iordane contra orientalem plagam 10:33 From Jordan eastward, all the land of Gilead, the

Gadites, and the Reubenites, and the Manassites, from Aroer, which is by the river Arnon, even Gilead and

omnem terram Galaad et Gad et Ruben et Manasse ab Aroer Bashan.

quae est super torrentem Arnon et Galaad et Basan 10:34 10:34 Now the rest of the acts of Jehu, and all that he did, and all his might, are they not written in the book of the chronicles of the kings of Israel?

reliqua autem uerborum Hieu et uniuersa quae fecit et fortitudo eius nonne haec scripta sunt in libro uerborum dierum regum Israhel 10:35 et dormiuit Hieu cum patribus suis se- 10:35 And Jehu slept with his fathers: and they buried him

in Samaria. And Jehoahaz his son reigned in his stead.

pelieruntque eum in Samaria et regnauit Ioachaz filius eius pro eo 10:36 dies autem quos regnauit Hieu super Israhel 10:36 And the time that Jehu reigned over Israel in

Samaria was twenty and eight years.

uiginti et octo anni sunt in Samaria

11:1 Athalia uero mater Ahaziae uidens mortuum filium 11:1 And when Athaliah the mother of Ahaziah saw that

her son was dead, she arose and destroyed all the seed royal.

suum surrexit et interfecit omne semen regium 11:2 tol-

11:2 But Jehosheba, the daughter of king Joram, sister of Ahaziah, took Joash the son of Ahaziah, and stole him from among the king’s sons which were slain; and they hid him, even him and his nurse, in the bedchamber from Athaliah, so that he was not slain.

lens autem Iosaba filia regis Ioram soror Ahaziae Ioas filium Ahaziae furata est eum de medio filiorum regis qui interficiebantur et nutricem eius de triclinio et abscondit eum a facie Athaliae ut non interficeretur 11:3 eratque cum ea in 11:3 And he was with her hid in the house of the LORD

six years. And Athaliah did reign over the land.

domo Domini clam sex annis porro Athalia regnauit super terram 11:4 anno autem septimo misit Ioiada et adsumens 11:4 And the seventh year Jehoiada sent and fetched the

rulers over hundreds, with the captains and the guard, and brought them to him into the house of the LORD, and made a covenant with them, and took an oath of them in the house of the LORD, and shewed them the king’s son.

centuriones et milites introduxit ad se in templum Domini pepigitque cum eis foedus et adiurans eos in domo Domini ostendit eis filium regis 11:5 et praecepit illis dicens iste 11:5 And he commanded them, saying, This is the thing

that ye shall do; A third part of you that enter in on the sabbath shall even be keepers of the watch of the king’s

sermo quem facere debetis 11:6 tertia pars uestrum introeat house;

11:6 And a third part shall be at the gate of Sur; and a third part at the gate behind the guard: so shall ye keep the watch of the house, that it be not broken down.

sabbato et obseruet excubitum domus regis tertia autem pars sit ad portam Sir et tertia pars ad portam quae est post habitaculum scutariorum et custodietis excubitum domus Messa

Biblia Sacra Vulgata

11:7 And two parts of all you that go forth on the sabbath, 11:7 duae uero partes e uobis omnes egredientes sabbato cu-

even they shall keep the watch of the house of the LORD about the king.

stodiant excubias domus Domini circum regem 11:8 et ual-

11:8 And ye shall compass the king round about, every man with his weapons in his hand: and he that cometh within the ranges, let him be slain: and be ye with the king as he goeth out and as he cometh in.

labitis eum habentes arma in manibus uestris si quis autem ingressus fuerit septum templi interficiatur eritisque cum

11:9 And the captains over the hundreds did according to rege introeunte et egrediente 11:9 et fecerunt centuriones

all things that Jehoiada the priest commanded: and they took every man his men that were to come in on the sabbath, with them that should go out on the sabbath, and came to Jehoiada the priest.

iuxta omnia quae praeceperat eis Ioiada sacerdos et adsumentes singuli uiros suos qui ingrediebantur sabbatum cum his qui egrediebantur e sabbato uenerunt ad Ioiada sacerdo-

11:10 And to the captains over hundreds did the priest give tem 11:10 qui dedit eis hastas et arma regis Dauid quae king David’s spears and shields, that were in the temple of the LORD.

erant in domo Domini 11:11 et steterunt singuli habentes

11:11 And the guard stood, every man with his weapons in his hand, round about the king, from the right corner of the temple to the left corner of the temple, along by the altar and the temple.

arma in manu sua a parte templi dextra usque ad partem

11:12 And he brought forth the king’s son, and put the sinistram altaris et aedis circum regem 11:12 produxitque

crown upon him, and gave him the testimony; and they made him king, and anointed him; and they clapped their

hands, and said, God save the king. filium regis et posuit super eum diadema et testimonium fe-

ceruntque eum regem et unxerunt et plaudentes manu dixe-

11:13 And when Athaliah heard the noise of the guard and runt uiuat rex 11:13 audiuit Athalia uocem currentis populi

of the people, she came to the people into the temple of the LORD.

et ingressa ad turbas in templum Domini 11:14 uidit regem

11:14 And when she looked, behold, the king stood by a pillar, as the manner was, and the princes and the trumpeters by the king, and all the people of the land rejoiced, and blew with trumpets: and Athaliah rent her clothes, and cried, Treason, Treason.

stantem super tribunal iuxta morem et cantores et tubas propter eum omnemque populum terrae laetantem et canentem tubis et scidit uestimenta sua clamauitque coniuratio coniu-

11:15 But Jehoiada the priest commanded the captains of ratio 11:15 praecepit autem Ioiada centurionibus qui erant

the hundreds, the officers of the host, and said unto them, Have her forth without the ranges: and him that followeth her kill with the sword. For the priest had said, Let her not be slain in the house of the LORD.

super exercitum et ait eis educite eam extra consepta templi et quicumque secutus eam fuerit feriatur gladio dixerat

11:16 And they laid hands on her; and she went by the enim sacerdos non occidatur in templo Domini 11:16 inpo-

way by the which the horses came into the king’s house: and there was she slain.

sueruntque ei manus et inpegerunt eam per uiam introitus

11:17 And Jehoiada made a covenant between the LORD equorum iuxta palatium et interfecta est ibi 11:17 pepigit

and the king and the people, that they should be the LORD’s people; between the king also and the people.

igitur Ioiada foedus inter Dominum et inter regem et inter populum ut esset populus Domini et inter regem et popu-

11:18 And all the people of the land went into the house lum 11:18 ingressusque est omnis populus terrae templum

of Baal, and brake it down; his altars and his images brake they in pieces thoroughly, and slew Mattan the priest of Baal before the altars. And the priest appointed officers over the house of the LORD.

Baal et destruxerunt aras eius et imagines contriuerunt ualide Matthan quoque sacerdotem Baal occiderunt coram altari et

11:19 And he took the rulers over hundreds, and the cap- posuit sacerdos custodias in domo Domini 11:19 tulitque

tains, and the guard, and all the people of the land; and they brought down the king from the house of the LORD, and came by the way of the gate of the guard to the king’s house. And he sat on the throne of the kings.

centuriones et Cherethi et Felethi legiones et omnem populum terrae deduxeruntque regem de domo Domini et uene-

Liber II Regum

runt per uiam portae scutariorum in palatium et sedit super thronum regum 11:20 laetatusque est omnis populus terrae 11:20 And all the people of the land rejoiced, and the city

was in quiet: and they slew Athaliah with the sword beside the king’s house.

et ciuitas conquieuit Athalia autem occisa est gladio in domo regis 11:21 septemque annorum erat Ioas cum regnare coe- 11:21 Seven years old was Jehoash when he began to

reign.

pisset

12:1 anno septimo Hieu regnauit Ioas quadraginta annis 12:1 In the seventh year of Jehu Jehoash began to reign;

and forty years reigned he in Jerusalem. And his mother’s name was Zibiah of Beersheba.

regnauit in Hierusalem nomen matris eius Sebia de Bersabee 12:2 fecitque Ioas rectum coram Domino cunctis die- 12:2 And Jehoash did that which was right in the sight

of the LORD all his days wherein Jehoiada the priest instructed him.

bus quibus docuit eum Ioiada sacerdos 12:3 uerumtamen

12:3 But the high places were not taken away: the people still sacrificed and burnt incense in the high places.

excelsa non abstulit adhuc populus immolabat et adolebat in excelsis incensum 12:4 dixitque Ioas ad sacerdotes om- 12:4 And Jehoash said to the priests, All the money of

the dedicated things that is brought into the house of the LORD, even the money of every one that passeth the account, the money that every man is set at, and all the money that cometh into any man’s heart to bring into the house of the LORD,

nem pecuniam sanctorum quae inlata fuerit in templum Domini a praetereuntibus quae offertur pro pretio animae et quam sponte et arbitrio cordis sui inferunt in templum Domini 12:5 accipiant illam sacerdotes iuxta ordinem suum et 12:5 Let the priests take it to them, every man of his ac-

quaintance: and let them repair the breaches of the house, wheresoever any breach shall be found.

instaurent sarta tecta domus si quid necessarium uiderint instauratione 12:6 igitur usque ad uicesimum tertium annum 12:6 But it was so, that in the three and twentieth year of

king Jehoash the priests had not repaired the breaches of the house.

regis Ioas non instaurauerunt sacerdotes sarta tecta templi 12:7 uocauitque rex Ioas Ioiada pontificem et sacerdotes di- 12:7 Then king Jehoash called for Jehoiada the priest, and

the other priests, and said unto them, Why repair ye not the breaches of the house? now therefore receive no more money of your acquaintance, but deliver it for the breaches of the house.

cens eis quare sarta tecta non instaurastis templi nolite ergo amplius accipere pecuniam iuxta ordinem uestrum sed ad instaurationem templi reddite eam 12:8 prohibitique sunt 12:8 And the priests consented to receive no more money

of the people, neither to repair the breaches of the house.

sacerdotes ultra accipere pecuniam a populo et instaurare sarta tecta domus 12:9 et tulit Ioiada pontifex gazofilacium 12:9 But Jehoiada the priest took a chest, and bored a hole

in the lid of it, and set it beside the altar, on the right side as one cometh into the house of the LORD: and the priests that kept the door put therein all the money that was brought into the house of the LORD.

unum aperuitque foramen desuper et posuit illud iuxta altare ad dexteram ingredientium domum Domini mittebantque in eo sacerdotes qui custodiebant ostia omnem pecuniam quae deferebatur ad templum Domini 12:10 cumque uiderent ni- 12:10 And it was so, when they saw that there was much

money in the chest, that the king’s scribe and the high priest came up, and they put up in bags, and told the money

miam pecuniam esse in gazofilacio ascendebat scriba regis that was found in the house of the LORD. et pontifex effundebantque et numerabant pecuniam quae in-

ueniebatur in domo Domini 12:11 et dabant eam iuxta nu- 12:11 And they gave the money, being told, into the hands of them that did the work, that had the oversight of the house of the LORD: and they laid it out to the carpenters

merum atque mensuram in manu eorum qui praeerant ce- and builders, that wrought upon the house of the LORD,

mentariis domus Domini qui inpendebant eam in fabris li-

Biblia Sacra Vulgata

gnorum et in cementariis his qui operabantur in domo Do-

12:12 And to masons, and hewers of stone, and to buy mini 12:12 et sarta tecta faciebant et in his qui caedebant

timber and hewed stone to repair the breaches of the house of the LORD, and for all that was laid out for the house to

repair it. saxa et ut emerent ligna et lapides qui excidebantur ita ut

impleretur instauratio domus Domini in uniuersis quae indi-

12:13 Howbeit there were not made for the house of the gebant expensa ad muniendam domum 12:13 uerumtamen LORD bowls of silver, snuffers, basons, trumpets, any vessels of gold, or vessels of silver, of the money that was

brought into the house of the LORD: non fiebant ex eadem pecunia hydriae templi Domini et fu-

scinulae et turibula et tubae omne uas aureum et argenteum

12:14 But they gave that to the workmen, and repaired de pecunia quae inferebatur in templum Domini 12:14 his

therewith the house of the LORD.

enim qui faciebant opus dabatur ut instauraretur templum

12:15 Moreover they reckoned not with the men, into Domini 12:15 et non fiebat ratio his hominibus qui acci-

whose hand they delivered the money to be bestowed on workmen: for they dealt faithfully.

piebant pecuniam ut distribuerent eam artificibus sed in fide

12:16 The trespass money and sin money was not brought tractabant eam 12:16 pecuniam uero pro delicto et pecu-

into the house of the LORD: it was the priests’.

niam pro peccatis non inferebant in templum Domini quia

12:17 Then Hazael king of Syria went up, and fought sacerdotum erat 12:17 tunc ascendit Azahel rex Syriae et

against Gath, and took it: and Hazael set his face to go up to Jerusalem.

pugnabat contra Geth cepitque eam et direxit faciem suam

12:18 And Jehoash king of Judah took all the hallowed ut ascenderet in Hierusalem 12:18 quam ob rem tulit Ioas

things that Jehoshaphat, and Jehoram, and Ahaziah, his fathers, kings of Judah, had dedicated, and his own hallowed things, and all the gold that was found in the treasures of the house of the LORD, and in the king’s house, and sent it to Hazael king of Syria: and he went away from Jerusalem.

rex Iuda omnia sanctificata quae consecrauerant Iosaphat et Ioram et Ahazia patres eius reges Iuda et quae ipse obtulerat et uniuersum argentum quod inueniri potuit in thesauris templi Domini et in palatio regis misitque Azaheli regi Syriae et

12:19 And the rest of the acts of Joash, and all that he did, recessit ab Hierusalem 12:19 reliqua autem sermonum Ioas

are they not written in the book of the chronicles of the kings of Judah?

et uniuersa quae fecit nonne haec scripta sunt in libro uer-

12:20 And his servants arose, and made a conspiracy, and borum dierum regum Iuda 12:20 surrexerunt autem serui slew Joash in the house of Millo, which goeth down to Silla.

eius et coniurauerunt inter se percusseruntque Ioas in domo

12:21 For Jozachar the son of Shimeath, and Jehozabad Mello in descensu Sela 12:21 Iozachar namque filius Se-

the son of Shomer, his servants, smote him, and he died; and they buried him with his fathers in the city of David:

and Amaziah his son reigned in his stead. math et Iozabad filius Somer serui eius percusserunt eum et mortuus est et sepelierunt eum cum patribus suis in ciuitate Dauid regnauitque Amasias filius eius pro eo

13:1 In the three and twentieth year of Joash the son of 13:1 anno uicesimo tertio Ioas filii Ahaziae regis Iudae re-

Ahaziah king of Judah Jehoahaz the son of Jehu began to reign over Israel in Samaria, and reigned seventeen years.

gnauit Ioachaz filius Hieu super Israhel in Samaria decem et

13:2 And he did that which was evil in the sight of the septem annis 13:2 et fecit malum coram Domino secutus-

LORD, and followed the sins of Jeroboam the son of Nebat, which made Israel to sin; he departed not there-

from. que est peccata Hieroboam filii Nabath qui peccare fecit Is-13:3 And the anger of the LORD was kindled against Is- rahel non declinauit ab eis 13:3 iratusque est furor Domini

rael, and he delivered them into the hand of Hazael king of Syria, and into the hand of Benhadad the son of Hazael, all their days.

Liber II Regum

contra Israhel et tradidit eos in manu Azahelis regis Syriae et in manu Benadad filii Azahel cunctis diebus 13:4 depreca- 13:4 And Jehoahaz besought the LORD, and the LORD

hearkened unto him: for he saw the oppression of Israel, because the king of Syria oppressed them.

tus est autem Ioachaz faciem Domini et audiuit eum Dominus uidit enim angustiam Israhel qua adtriuerat eos rex Syriae 13:5 et dedit Dominus Israheli saluatorem et liberatus 13:5 (And the LORD gave Israel a saviour, so that they

went out from under the hand of the Syrians: and the children of Israel dwelt in their tents, as beforetime.

est de manu Syriae habitaueruntque filii Israhel in tabernaculis suis sicut heri et nudius tertius 13:6 uerumtamen non 13:6 Nevertheless they departed not from the sins of the

house of Jeroboam, who made Israel sin, but walked therein: and there remained the grove also in Samaria.)

recesserunt a peccatis domus Hieroboam qui peccare fecit Israhel in ipsis ambulauerunt siquidem et lucus permansit in Samaria 13:7 et non sunt derelicti Ioachaz de populo nisi 13:7 Neither did he leave of the people to Jehoahaz but

fifty horsemen, and ten chariots, and ten thousand footmen; for the king of Syria had destroyed them, and had

quinquaginta equites et decem currus et decem milia pedi- made them like the dust by threshing. tum interfecerat enim eos rex Syriae et redeeerat quasi puluerem in tritura areae 13:8 reliqua autem sermonum Ioa- 13:8 Now the rest of the acts of Jehoahaz, and all that he

did, and his might, are they not written in the book of the chronicles of the kings of Israel?

chaz et uniuersa quae fecit sed et fortitudo eius nonne haec scripta sunt in libro sermonum dierum regum Israhel 13:9 13:9 And Jehoahaz slept with his fathers; and they buried

him in Samaria: and Joash his son reigned in his stead.

dormiuitque Ioachaz cum patribus suis et sepelierunt eum in Samaria regnauitque Ioas filius eius pro eo 13:10 anno tri- 13:10 In the thirty and seventh year of Joash king of Judah

began Jehoash the son of Jehoahaz to reign over Israel in Samaria, and reigned sixteen years.

cesimo septimo Ioas regis Iuda regnauit Ioas filius Ioachaz super Israhel in Samaria sedecim annis 13:11 et fecit quod 13:11 And he did that which was evil in the sight of the

LORD; he departed not from all the sins of Jeroboam the son of Nebat, who made Israel sin: but he walked therein.

malum est in conspectu Domini non declinauit ab omnibus peccatis Hieroboam filii Nabath qui peccare fecit Israhel in ipsis ambulauit 13:12 reliqua autem sermonum Ioas et uni- 13:12 And the rest of the acts of Joash, and all that he did,

and his might wherewith he fought against Amaziah king of Judah, are they not written in the book of the chronicles

uersa quae fecit sed et fortitudo eius quomodo pugnauerit of the kings of Israel? contra Amasiam regem Iuda nonne haec scripta sunt in libro

sermonum regum Israhel 13:13 et dormiuit Ioas cum pat- 13:13 And Joash slept with his fathers; and Jeroboam sat upon his throne: and Joash was buried in Samaria with the kings of Israel.

ribus suis Hieroboam autem sedit super solium eius porro Ioas sepultus est in Samaria cum regibus Israhel 13:14 He- 13:14 Now Elisha was fallen sick of his sickness whereof

he died. And Joash the king of Israel came down unto him, and wept over his face, and said, O my father, my father,

liseus autem aegrotabat infirmitate qua et mortuus est de- the chariot of Israel, and the horsemen thereof. scenditque ad eum Ioas rex Israhel et flebat coram eo dicebatque pater mi pater mi currus Israhel et auriga eius 13:15 13:15 And Elisha said unto him, Take bow and arrows.

And he took unto him bow and arrows.

et ait illi Heliseus adfer arcum et sagittas cumque adtulisset ad eum arcum et sagittas 13:16 dixit ad regem Israhel pone 13:16 And he said to the king of Israel, Put thine hand

upon the bow. And he put his hand upon it: and Elisha put his hands upon the king’s hands.

manum tuam super arcum et cum posuisset ille manum suam

Biblia Sacra Vulgata

13:17 And he said, Open the window eastward. And he superposuit Heliseus manus suas manibus regis 13:17 et ait

opened it. Then Elisha said, Shoot. And he shot. And he said, The arrow of the LORD’s deliverance, and the arrow of deliverance from Syria: for thou shalt smite the Syrians in Aphek, till thou have consumed them.

aperi fenestram orientalem cumque aperuisset dixit Heliseus iace sagittam et iecit et ait Heliseus sagitta salutis Domini et sagitta salutis contra Syriam percutiesque Syriam in Afec

13:18 And he said, Take the arrows. And he took them. donec consumas eam 13:18 et ait tolle sagittas qui cum tu-

And he said unto the king of Israel, Smite upon the ground. And he smote thrice, and stayed.

lisset rursum dixit ei percute iaculo terram et cum percus-

13:19 And the man of God was wroth with him, and said, sisset tribus uicibus et stetisset 13:19 iratus est contra eum

Thou shouldest have smitten five or six times; then hadst thou smitten Syria till thou hadst consumed it: whereas

now thou shalt smite Syria but thrice. uir Dei et ait si percussisses quinquies aut sexies siue septies

percussisses Syriam usque ad consummationem nunc autem

13:20 And Elisha died, and they buried him. And the tribus uicibus percuties eam 13:20 mortuus est ergo Heli-

bands of the Moabites invaded the land at the coming in of the year.

seus et sepelierunt eum latrunculi quoque de Moab uenerunt

13:21 And it came to pass, as they were burying a man, in terra in ipso anno 13:21 quidam autem sepelientes homi-

that, behold, they spied a band of men; and they cast the man into the sepulchre of Elisha: and when the man was let down, and touched the bones of Elisha, he revived, and stood up on his feet.

nem uiderunt latrunculos et proiecerunt cadauer in sepulchro Helisei quod ambulauit et tetigit ossa Helisei et reuixit homo

13:22 But Hazael king of Syria oppressed Israel all the et stetit super pedes suos 13:22 igitur Azahel rex Syriae ad-

days of Jehoahaz.

13:23 And the LORD was gracious unto them, and had flixit Israhel cunctis diebus Ioachaz 13:23 et misertus est

compassion on them, and had respect unto them, because of his covenant with Abraham, Isaac, and Jacob, and would not destroy them, neither cast he them from his presence as yet.

Dominus eorum et reuersus est ad eos propter pactum suum quod habebat cum Abraham Isaac et Iacob et noluit disperdere eos neque proicere penitus usque in praesens tempus

13:24 So Hazael king of Syria died; and Benhadad his son 13:24 mortuus est autem Azahel rex Syriae et regnauit Be-

reigned in his stead.

13:25 And Jehoash the son of Jehoahaz took again out of nadad filius eius pro eo 13:25 porro Ioas filius Ioachaz tulit

the hand of Benhadad the son of Hazael the cities, which he had taken out of the hand of Jehoahaz his father by war. Three times did Joash beat him, and recovered the cities of Israel.

urbes de manu Benadad filii Azahel quas tulerat de manu Ioachaz patris sui iure proelii tribus uicibus percussit eum Ioas et reddidit ciuitates Israheli

14:1 In the second year of Joash son of Jehoahaz king of 14:1 anno secundo Ioas filii Ioachaz regis Israhel regna-

Israel reigned Amaziah the son of Joash king of Judah.

14:2 He was twenty and five years old when he began to uit Amasias filius Ioas regis Iuda 14:2 uiginti quinque an-

reign, and reigned twenty and nine years in Jerusalem. And his mother’s name was Jehoaddan of Jerusalem.

norum erat cum regnare coepisset uiginti autem et nouem annis regnauit in Hierusalem nomen matris eius Ioaden de

14:3 And he did that which was right in the sight of the Hierusalem 14:3 et fecit rectum coram Domino uerumta-

LORD, yet not like David his father: he did according to all things as Joash his father did.

men non ut Dauid pater eius iuxta omnia quae fecit Ioas pa-

14:4 Howbeit the high places were not taken away: as ter suus fecit 14:4 nisi hoc tantum quod excelsa non abs-

yet the people did sacrifice and burnt incense on the high places.

tulit adhuc enim populus immolabat et adolebat in excelsis

14:5 And it came to pass, as soon as the kingdom was 14:5 cumque obtinuisset regnum percussit seruos suos qui

confirmed in his hand, that he slew his servants which had slain the king his father.

Liber II Regum

interfecerant regem patrem suum 14:6 filios autem eorum 14:6 But the children of the murderers he slew not: ac-

cording unto that which is written in the book of the law of Moses, wherein the LORD commanded, saying, The fathers shall not be put to death for the children, nor the children be put to death for the fathers; but every man shall be put to death for his own sin.

qui occiderant non occidit iuxta quod scriptum est in libro legis Mosi sicut praecepit Dominus dicens non morientur patres pro filiis neque filii morientur pro patribus sed unusquisque in peccato suo morietur 14:7 ipse percussit Edom 14:7 He slew of Edom in the valley of salt ten thousand,

and took Selah by war, and called the name of it Joktheel unto this day.

in ualle Salinarum decem milia et adprehendit Petram in proelio uocauitque nomen eius Iecethel usque in praesentem diem 14:8 tunc misit Amasias nuntios ad Ioas filium 14:8 Then Amaziah sent messengers to Jehoash, the son

of Jehoahaz son of Jehu, king of Israel, saying, Come, let us look one another in the face.

Ioachaz filii Hieu regis Israhel dicens ueni et uideamus nos 14:9 remisitque Ioas rex Israhel ad Amasiam regem Iuda di- 14:9 And Jehoash the king of Israel sent to Amaziah king

of Judah, saying, The thistle that was in Lebanon sent to the cedar that was in Lebanon, saying, Give thy daughter to my son to wife: and there passed by a wild beast that was in Lebanon, and trode down the thistle.

cens carduus Libani misit ad cedrum quae est in Libano dicens da filiam tuam filio meo uxorem transieruntque bestiae saltus quae sunt in Libano et conculcauerunt carduum 14:10 percutiens inualuisti super Edom et subleuauit te cor 14:10 Thou hast indeed smitten Edom, and thine heart

hath lifted thee up: glory of this, and tarry at home: for why shouldest thou meddle to thy hurt, that thou shouldest

tuum contentus esto gloria et sede in domo tua quare prouo- fall, even thou, and Judah with thee? cas malum ut cadas tu et Iuda tecum 14:11 et non adquie- 14:11 But Amaziah would not hear. Therefore Jehoash

king of Israel went up; and he and Amaziah king of Judah looked one another in the face at Bethshemesh, which

uit Amasias ascenditque Ioas rex Israhel et uiderunt se ipse belongeth to Judah. et Amasias rex Iuda in Bethsames oppido Iudae 14:12 per- 14:12 And Judah was put to the worse before Israel; and

they fled every man to their tents.

cussusque est Iuda coram Israhel et fugerunt unusquisque in tabernacula sua 14:13 Amasiam uero regem Iuda filium 14:13 And Jehoash king of Israel took Amaziah king

of Judah, the son of Jehoash the son of Ahaziah, at Bethshemesh, and came to Jerusalem, and brake down the wall of Jerusalem from the gate of Ephraim unto the corner gate, four hundred cubits.

Ioas filii Ahaziae cepit Ioas rex Israhel in Bethsames et adduxit eum in Hierusalem et interrupit murum Hierusalem a porta Ephraim usque ad portam Anguli quadringentis cubitis 14:14 tulitque omne aurum et argentum et uniuersa uasa quae 14:14 And he took all the gold and silver, and all the ves-

sels that were found in the house of the LORD, and in the treasures of the king’s house, and hostages, and returned

inuenta sunt in domo Domini et in thesauris regis et obsi- to Samaria. des et reuersus est Samariam 14:15 reliqua autem uerborum 14:15 Now the rest of the acts of Jehoash which he did,

and his might, and how he fought with Amaziah king of Judah, are they not written in the book of the chronicles of

Ioas quae fecit et fortitudo eius qua pugnauit contra Ama- the kings of Israel? siam regem Iuda nonne haec scripta sunt in libro sermonum dierum regum Israhel 14:16 dormiuitque Ioas cum patribus 14:16 And Jehoash slept with his fathers, and was buried

in Samaria with the kings of Israel; and Jeroboam his son reigned in his stead.

suis et sepultus est in Samaria cum regibus Israhel et regnauit Hieroboam filius eius pro eo 14:17 uixit autem Amasias 14:17 And Amaziah the son of Joash king of Judah lived

after the death of Jehoash son of Jehoahaz king of Israel fifteen years.

filius Ioas rex Iuda postquam mortuus est Ioas filius Ioachaz regis Israhel uiginti quinque annis 14:18 reliqua autem ser- 14:18 And the rest of the acts of Amaziah, are they not

written in the book of the chronicles of the kings of Judah?

Biblia Sacra Vulgata

monum Amasiae nonne haec scripta sunt in libro sermonum

14:19 Now they made a conspiracy against him in dierum regum Iuda 14:19 factaque est contra eum coniu-

Jerusalem: and he fled to Lachish; but they sent after him to Lachish, and slew him there.

ratio in Hierusalem at ille fugit in Lachis miseruntque post

14:20 And they brought him on horses: and he was buried eum in Lachis et interfecerunt eum ibi 14:20 et asportaue-

at Jerusalem with his fathers in the city of David.

runt in equis sepultusque est in Hierusalem cum patribus suis

14:21 And all the people of Judah took Azariah, which in ciuitate Dauid 14:21 tulit autem uniuersus populus Iudae

was sixteen years old, and made him king instead of his father Amaziah.

Azariam annos natum sedecim et constituerunt eum regem

14:22 He built Elath, and restored it to Judah, after that the pro patre eius Amasia 14:22 ipse aedificauit Ahilam et reking slept with his fathers.

stituit eam Iudae postquam dormiuit rex cum patribus suis

14:23 In the fifteenth year of Amaziah the son of Joash 14:23 anno quintodecimo Amasiae filii Ioas regis Iuda reg-

king of Judah Jeroboam the son of Joash king of Israel began to reign in Samaria, and reigned forty and one years.

nauit Hieroboam filius Ioas regis Israhel in Samaria quad-

14:24 And he did that which was evil in the sight of the raginta et uno anno 14:24 et fecit quod malum est coram

LORD: he departed not from all the sins of Jeroboam the son of Nebat, who made Israel to sin.

Domino non recessit ab omnibus peccatis Hieroboam filii

14:25 He restored the coast of Israel from the entering of Nabath qui peccare fecit Israhel 14:25 ipse restituit termi-

Hamath unto the sea of the plain, according to the word of the LORD God of Israel, which he spake by the hand of his servant Jonah, the son of Amittai, the prophet, which was of Gathhepher.

nos Israhel ab introitu Emath usque ad mare Solitudinis iuxta sermonem Domini Dei Israhel quem locutus est per seruum suum Ionam filium Amathi prophetam qui erat de Geth quae

14:26 For the LORD saw the affliction of Israel, that it was est in Opher 14:26 uidit enim Dominus adflictionem Isravery bitter: for there was not any shut up, nor any left, nor any helper for Israel.

hel amaram nimis et quod consumpti essent usque ad clausos carcere et extremos et non esset qui auxiliaretur Israhel

14:27 And the LORD said not that he would blot out the 14:27 nec locutus est Dominus ut deleret nomen Israhel sub

name of Israel from under heaven: but he saved them by the hand of Jeroboam the son of Joash.

caelo sed saluauit eos in manu Hieroboam filii Ioas 14:28

14:28 Now the rest of the acts of Jeroboam, and all that he did, and his might, how he warred, and how he recovered Damascus, and Hamath, which belonged to Judah, for Israel, are they not written in the book of the chronicles of the kings of Israel?

reliqua autem sermonum Hieroboam et uniuersa quae fecit et fortitudo eius qua proeliatus est et quomodo restituit Damascum et Emath Iudae in Israhel nonne haec scripta sunt

14:29 And Jeroboam slept with his fathers, even with the in libro sermonum dierum regum Israhel 14:29 dormiuitkings of Israel; and Zachariah his son reigned in his stead.

que Hieroboam cum patribus suis regibus Israhel et regnauit Zaccharias filius eius pro eo

15:1 In the twenty and seventh year of Jeroboam king of 15:1 anno uicesimo septimo Hieroboam regis Israhel reg-

Israel began Azariah son of Amaziah king of Judah to reign.

nauit Azarias filius Amasiae regis Iudae 15:2 sedecim an-

15:2 Sixteen years old was he when he began to reign, and he reigned two and fifty years in Jerusalem. And his mother’s name was Jecholiah of Jerusalem.

norum erat cum regnare coepisset et quinquaginta duobus annis regnauit in Hierusalem nomen matris eius Iecelia de

15:3 And he did that which was right in the sight of the Hierusalem 15:3 fecitque quod erat placitum coram Do-

LORD, according to all that his father Amaziah had done;

Liber II Regum

mino iuxta omnia quae fecit Amasias pater eius 15:4 ue- 15:4 Save that the high places were not removed: the peo-

ple sacrificed and burnt incense still on the high places.

rumtamen excelsa non est demolitus adhuc populus sacrificabat et adolebat incensum in excelsis 15:5 percussit autem 15:5 And the LORD smote the king, so that he was a leper

unto the day of his death, and dwelt in a several house. And Jotham the king’s son was over the house, judging

Dominus regem et fuit leprosus usque in diem mortis suae the people of the land. et habitabat in domo libera seorsum Ioatham uero filius regis gubernabat palatium et iudicabat populum terrae 15:6 reli- 15:6 And the rest of the acts of Azariah, and all that he

did, are they not written in the book of the chronicles of the kings of Judah?

qua autem sermonum Azariae et uniuersa quae fecit nonne haec scripta sunt in libro uerborum dierum regum Iuda 15:7 15:7 So Azariah slept with his fathers; and they buried him

with his fathers in the city of David: and Jotham his son reigned in his stead.

et dormiuit Azarias cum patribus suis sepelieruntque eum cum maioribus suis in ciuitate Dauid et regnauit Ioatham filius eius pro eo 15:8 anno tricesimo octauo Azariae re- 15:8 In the thirty and eighth year of Azariah king of Ju-

dah did Zachariah the son of Jeroboam reign over Israel in Samaria six months.

gis Iudae regnauit Zaccharias filius Hieroboam super Israhel in Samaria sex mensibus 15:9 et fecit quod malum est 15:9 And he did that which was evil in the sight of the

LORD, as his fathers had done: he departed not from the sins of Jeroboam the son of Nebat, who made Israel to sin.

coram Domino sicut fecerant patres eius non recessit a peccatis Hieroboam filii Nabath qui peccare fecit Israhel 15:10 15:10 And Shallum the son of Jabesh conspired against

him, and smote him before the people, and slew him, and reigned in his stead.

coniurauit autem contra eum Sellum filius Iabes percussitque eum palam et interfecit regnauitque pro eo 15:11 reli- 15:11 And the rest of the acts of Zachariah, behold, they

are written in the book of the chronicles of the kings of Israel.

qua autem uerborum Zacchariae nonne haec scripta sunt in libro sermonum dierum regum Israhel 15:12 ipse est sermo 15:12 This was the word of the LORD which he spake

unto Jehu, saying, Thy sons shall sit on the throne of Israel unto the fourth generation. And so it came to pass.

Domini quem locutus est ad Hieu dicens filii usque ad quartam generationem sedebunt de te super thronum Israhel factumque est ita 15:13 Sellum filius Iabes regnauit tricesimo 15:13 Shallum the son of Jabesh began to reign in the nine

and thirtieth year of Uzziah king of Judah; and he reigned a full month in Samaria.

nono anno Azariae regis Iudae regnauit autem uno mense in Samaria 15:14 et ascendit Manahem filius Gaddi de Thersa 15:14 For Menahem the son of Gadi went up from Tirzah,

and came to Samaria, and smote Shallum the son of Jabesh in Samaria, and slew him, and reigned in his stead.

uenitque Samariam et percussit Sellum filium Iabes in Samaria et interfecit eum regnauitque pro eo 15:15 reliqua au- 15:15 And the rest of the acts of Shallum, and his conspir-

acy which he made, behold, they are written in the book of the chronicles of the kings of Israel.

tem uerborum Sellum et coniuratio eius per quam tetendit insidias nonne haec scripta sunt in libro sermonum dierum regum Israhel 15:16 tunc percussit Manahem Thapsam et 15:16 Then Menahem smote Tiphsah, and all that were

therein, and the coasts thereof from Tirzah: because they opened not to him, therefore he smote it; and all the

omnes qui erant in ea et terminos eius de Thersa noluerant women therein that were with child he ripped up. enim aperire ei et interfecit omnes praegnantes eius et scidit eas 15:17 anno tricesimo nono Azariae regis Iuda regnauit 15:17 In the nine and thirtieth year of Azariah king of Ju-

dah began Menahem the son of Gadi to reign over Israel, and reigned ten years in Samaria.

Manahem filius Gaddi super Israhel decem annis in Samaria

Biblia Sacra Vulgata

15:18 And he did that which was evil in the sight of the 15:18 fecitque quod erat malum coram Domino non recessit

LORD: he departed not all his days from the sins of Jeroboam the son of Nebat, who made Israel to sin.

a peccatis Hieroboam filii Nabath qui peccare fecit Israhel

15:19 And Pul the king of Assyria came against the land: cunctis diebus eius 15:19 ueniebat Phul rex Assyriorum in

and Menahem gave Pul a thousand talents of silver, that his hand might be with him to confirm the kingdom in his

hand. terram et dabat Manahem Phul mille talenta argenti ut esset 15:20 And Menahem exacted the money of Israel, even of ei in auxilio et firmaret regnum eius 15:20 indixitque Ma-

all the mighty men of wealth, of each man fifty shekels of silver, to give to the king of Assyria. So the king of

Assyria turned back, and stayed not there in the land. nahem argentum super Israhel cunctis potentibus et diuitibus ut daret regi Assyriorum quinquaginta siclos argenti per singulos reuersusque est rex Assyriorum et non est moratus in

15:21 And the rest of the acts of Menahem, and all that he terra 15:21 reliqua autem sermonum Manahem et uniuersa

did, are they not written in the book of the chronicles of the kings of Israel?

quae fecit nonne haec scripta sunt in libro sermonum die-

15:22 And Menahem slept with his fathers; and Pekahiah rum regum Israhel 15:22 et dormiuit Manahem cum patrihis son reigned in his stead.

15:23 In the fiftieth year of Azariah king of Judah bus suis regnauitque Phaceia filius eius pro eo 15:23 anno

Pekahiah the son of Menahem began to reign over Israel in Samaria, and reigned two years.

quinquagesimo Azariae regis Iudae regnauit Phaceia filius

15:24 And he did that which was evil in the sight of the Manahem super Israhel in Samaria biennio 15:24 et fecit

LORD: he departed not from the sins of Jeroboam the son of Nebat, who made Israel to sin.

quod erat malum coram Domino non recessit a peccatis Hie-

15:25 But Pekah the son of Remaliah, a captain of his, roboam filii Nabath qui peccare fecit Israhel 15:25 coniura-

conspired against him, and smote him in Samaria, in the palace of the king’s house, with Argob and Arieh, and with him fifty men of the Gileadites: and he killed him, and reigned in his room.

uit autem aduersum eum Phacee filius Romeliae dux eius et percussit eum in Samaria in turre domus regiae iuxta Argob et iuxta Ari et cum eo quinquaginta uiros de filiis Galaa-

15:26 And the rest of the acts of Pekahiah, and all that he ditarum et interfecit eum regnauitque pro eo 15:26 reliqua

did, behold, they are written in the book of the chronicles of the kings of Israel.

autem sermonum Phaceia et uniuersa quae fecit nonne haec

15:27 In the two and fiftieth year of Azariah king of Judah scripta sunt in libro sermonum dierum regum Israhel 15:27

Pekah the son of Remaliah began to reign over Israel in Samaria, and reigned twenty years.

anno quinquagesimo secundo Azariae regis Iudae regnauit Phacee filius Romeliae super Israhel in Samaria uiginti an-

15:28 And he did that which was evil in the sight of the nis 15:28 et fecit quod malum erat coram Domino non re-

LORD: he departed not from the sins of Jeroboam the son of Nebat, who made Israel to sin.

cessit a peccatis Hieroboam filii Nabath qui peccare fecit

15:29 In the days of Pekah king of Israel came Tiglath- Israhel 15:29 in diebus Phacee regis Israhel uenit Theglath-

pileser king of Assyria, and took Ijon, and Abelbethmaachah, and Janoah, and Kedesh, and Hazor, and Gilead, and Galilee, all the land of Naphtali, and carried them captive to Assyria.

falassar rex Assur et cepit Aiom et Abel domum Maacha et Ianoe et Cedes et Asor et Galaad et Galileam uniuersam ter-

15:30 And Hoshea the son of Elah made a conspiracy ram Nepthalim et transtulit eos in Assyrios 15:30 coniurauit

against Pekah the son of Remaliah, and smote him, and slew him, and reigned in his stead, in the twentieth year of

Jotham the son of Uzziah. autem et tetendit insidias Osee filius Hela contra Phacee fi-

lium Romeliae et percussit eum et interfecit regnauitque pro

15:31 And the rest of the acts of Pekah, and all that he did, eo uicesimo anno Ioatham filii Oziae 15:31 reliqua autem behold, they are written in the book of the chronicles of the kings of Israel.

Liber II Regum

sermonum Phacee et uniuersa quae fecit nonne haec scripta sunt in libro sermonum dierum regum Israhel 15:32 anno 15:32 In the second year of Pekah the son of Remaliah

king of Israel began Jotham the son of Uzziah king of Judah to reign.

secundo Phacee filii Romeliae regis Israhel regnauit Ioatham filius Oziae regis Iuda 15:33 uiginti quinque annorum erat 15:33 Five and twenty years old was he when he began to

reign, and he reigned sixteen years in Jerusalem. And his mother’s name was Jerusha, the daughter of Zadok.

cum regnare coepisset et sedecim annis regnauit in Hierusalem nomen matris eius Hierusa filia Sadoc 15:34 fecitque 15:34 And he did that which was right in the sight of the

LORD: he did according to all that his father Uzziah had done.

quod erat placitum coram Domino iuxta omnia quae fecerat Ozias pater suus operatus est 15:35 uerumtamen excelsa 15:35 Howbeit the high places were not removed: the peo-

ple sacrificed and burned incense still in the high places. He built the higher gate of the house of the LORD.

non abstulit adhuc populus immolabat et adolebat incensum in excelsis ipse aedificauit portam domus Domini sublimissimam 15:36 reliqua autem sermonum Ioatham et uniuersa 15:36 Now the rest of the acts of Jotham, and all that he

did, are they not written in the book of the chronicles of the kings of Judah?

quae fecit nonne haec scripta sunt in libro uerborum dierum regum Iuda 15:37 in diebus illis coepit Dominus mittere in 15:37 In those days the LORD began to send against Judah

Rezin the king of Syria, and Pekah the son of Remaliah.

Iudam Rasin regem Syriae et Phacee filium Romeliae 15:38 15:38 And Jotham slept with his fathers, and was buried

with his fathers in the city of David his father: and Ahaz his son reigned in his stead.

et dormiuit Ioatham cum patribus suis sepultusque est cum eis in ciuitate Dauid patris sui et regnauit Ahaz filius eius pro eo

16:1 anno septimodecimo Phacee filii Romeliae regnauit 16:1 In the seventeenth year of Pekah the son of Remaliah

Ahaz the son of Jotham king of Judah began to reign.

Ahaz filius Ioatham regis Iuda 16:2 uiginti annorum erat 16:2 Twenty years old was Ahaz when he began to reign,

and reigned sixteen years in Jerusalem, and did not that which was right in the sight of the LORD his God, like

Ahaz cum regnare coepisset et sedecim annis regnauit in David his father. Hierusalem non fecit quod erat placitum in conspectu Domini Dei sui sicut Dauid pater eius 16:3 sed ambulauit in 16:3 But he walked in the way of the kings of Israel, yea,

and made his son to pass through the fire, according to the abominations of the heathen, whom the LORD cast out

uia regum Israhel insuper et filium suum consecrauit transfe- from before the children of Israel. rens per ignem secundum idola gentium quae dissipauit Dominus coram filiis Israhel 16:4 immolabat quoque uictimas 16:4 And he sacrificed and burnt incense in the high

places, and on the hills, and under every green tree.

et adolebat incensum in excelsis et in collibus et sub omni ligno frondoso 16:5 tunc ascendit Rasin rex Syriae et Pha- 16:5 Then Rezin king of Syria and Pekah son of Remaliah

king of Israel came up to Jerusalem to war: and they besieged Ahaz, but could not overcome him.

cee filius Romeliae rex Israhel in Hierusalem ad proeliandum cumque obsiderent Ahaz non ualuerunt superare eum 16:6 in tempore illo restituit Rasin rex Syriae Ahilam Syriae 16:6 At that time Rezin king of Syria recovered Elath to

Syria, and drave the Jews from Elath: and the Syrians came to Elath, and dwelt there unto this day.

et eiecit Iudaeos de Ahilam et Idumei uenerunt in Ahilam et habitauerunt ibi usque in diem hanc 16:7 misit autem Ahaz 16:7 So Ahaz sent messengers to Tiglathpileser king of

Assyria, saying, I am thy servant and thy son: come up, and save me out of the hand of the king of Syria, and out

nuntios ad Theglathfalassar regem Assyriorum dicens ser- of the hand of the king of Israel, which rise up against me.

Biblia Sacra Vulgata

uus tuus et filius tuus ego sum ascende et saluum me fac de manu regis Syriae et de manu regis Israhel qui consur-

16:8 And Ahaz took the silver and gold that was found in rexerunt aduersum me 16:8 et cum collegisset argentum et

the house of the LORD, and in the treasures of the king’s house, and sent it for a present to the king of Assyria.

aurum quod inuenire potuit in domo Domini et in thesauris

16:9 And the king of Assyria hearkened unto him: for the regis misit regi Assyriorum munera 16:9 qui et adquieuit

king of Assyria went up against Damascus, and took it, and carried the people of it captive to Kir, and slew Rezin.

uoluntati eius ascendit enim rex Assyriorum in Damascum et uastauit eam et transtulit habitatores eius Cyrenen Rasin au-

16:10 And king Ahaz went to Damascus to meet Tiglath- tem interfecit 16:10 perrexitque rex Ahaz in occursum The-

pileser king of Assyria, and saw an altar that was at Damascus: and king Ahaz sent to Urijah the priest the fashion of the altar, and the pattern of it, according to all the workmanship thereof.

glathfalassar regis Assyriorum in Damascum cumque uidisset altare Damasci misit rex Ahaz ad Vriam sacerdotem ex-

16:11 And Urijah the priest built an altar according to all emplar eius et similitudinem iuxta omne opus eius 16:11 that king Ahaz had sent from Damascus: so Urijah the priest made it against king Ahaz came from Damascus.

extruxitque Vrias sacerdos altare iuxta omnia quae praeceperat rex Ahaz de Damasco ita fecit Vrias sacerdos donec

16:12 And when the king was come from Damascus, the ueniret rex Ahaz de Damasco 16:12 cumque uenisset rex

king saw the altar: and the king approached to the altar, and offered thereon.

de Damasco uidit altare et ueneratus est illud ascenditque et

16:13 And he burnt his burnt offering and his meat offer- immolauit holocausta et sacrificium suum 16:13 et libauit

ing, and poured his drink offering, and sprinkled the blood of his peace offerings, upon the altar.

libamina et fudit sanguinem pacificorum quae obtulerat su-

16:14 And he brought also the brasen altar, which was be- per altare 16:14 porro altare aeneum quod erat coram Dofore the LORD, from the forefront of the house, from between the altar and the house of the LORD, and put it on

the north side of the altar. mino transtulit de facie templi et de loco altaris et de loco

templi Domini posuitque illud ex latere altaris ad aquilonem

16:15 And king Ahaz commanded Urijah the priest, say- 16:15 praecepit quoque rex Ahaz Vriae sacerdoti dicens su-

ing, Upon the great altar burn the morning burnt offering, and the evening meat offering, and the king’s burnt sacrifice, and his meat offering, with the burnt offering of all the people of the land, and their meat offering, and their drink offerings; and sprinkle upon it all the blood of the burnt offering, and all the blood of the sacrifice: and the brasen altar shall be for me to enquire by.

per altare maius offer holocaustum matutinum et sacrificium uespertinum et holocaustum regis et sacrificium eius et holocaustum uniuersi populi terrae et sacrificia eorum et libamina eorum et omnem sanguinem holocausti et uniuersum sanguinem uictimae super illud effundes altare uero aeneum

16:16 Thus did Urijah the priest, according to all that king erit paratum ad uoluntatem meam 16:16 fecit igitur Vrias Ahaz commanded.

16:17 And king Ahaz cut off the borders of the bases, and sacerdos iuxta omnia quae praeceperat rex Ahaz 16:17 tu-

removed the laver from off them; and took down the sea from off the brasen oxen that were under it, and put it upon

the pavement of stones. lit autem rex Ahaz celatas bases et luterem qui erat desu-

per et mare deposuit de bubus aeneis qui sustentabant illud

16:18 And the covert for the sabbath that they had built in et posuit super pauimentum stratum lapide 16:18 Musach the house, and the king’s entry without, turned he from the house of the LORD for the king of Assyria.

quoque sabbati quod aedificauerat in templo et ingressum regis exterius conuertit in templo Domini propter regem As-

Liber II Regum

syriorum 16:19 reliqua autem uerborum Ahaz quae fecit 16:19 Now the rest of the acts of Ahaz which he did, are

they not written in the book of the chronicles of the kings of Judah?

nonne haec scripta sunt in libro sermonum dierum regum Iuda 16:20 dormiuitque Ahaz cum patribus suis et sepultus 16:20 And Ahaz slept with his fathers, and was buried

with his fathers in the city of David: and Hezekiah his son reigned in his stead.

est cum eis in ciuitate Dauid et regnauit Ezechias filius eius pro eo

17:1 anno duodecimo Ahaz regis Iuda regnauit Osee filius 17:1 In the twelfth year of Ahaz king of Judah began

Hoshea the son of Elah to reign in Samaria over Israel nine years.

Hela in Samaria super Israhel nouem annis 17:2 fecitque

17:2 And he did that which was evil in the sight of the LORD, but not as the kings of Israel that were before him.

malum coram Domino sed non sicut reges Israhel qui ante eum fuerant 17:3 contra hunc ascendit Salmanassar rex As- 17:3 Against him came up Shalmaneser king of Assyria;

and Hoshea became his servant, and gave him presents.

syriorum et factus est ei Osee seruus reddebatque illi tributa 17:4 cumque deprehendisset rex Assyriorum Osee quod re- 17:4 And the king of Assyria found conspiracy in Hoshea:

for he had sent messengers to So king of Egypt, and brought no present to the king of Assyria, as he had done year by year: therefore the king of Assyria shut him up, and bound him in prison.

bellare nitens misisset nuntios ad Sua regem Aegypti ne praestaret tributa regi Assyriorum sicut singulis annis solitus erat obsedit eum et uinctum misit in carcerem 17:5 perua- 17:5 Then the king of Assyria came up throughout all the

land, and went up to Samaria, and besieged it three years.

gatusque est omnem terram et ascendens Samariam obsedit eam tribus annis 17:6 anno autem nono Osee cepit rex As- 17:6 In the ninth year of Hoshea the king of Assyria took

Samaria, and carried Israel away into Assyria, and placed them in Halah and in Habor by the river of Gozan, and in

syriorum Samariam et transtulit Israhel in Assyrios posuit- the cities of the Medes. que eos in Ala et in Habor iuxta fluuium Gozan in ciuitatibus Medorum 17:7 factum est enim cum peccassent filii Israhel 17:7 For so it was, that the children of Israel had sinned

against the LORD their God, which had brought them up out of the land of Egypt, from under the hand of Pharaoh

Domino Deo suo qui eduxerat eos de terra Aegypti de manu king of Egypt, and had feared other gods, Pharaonis regis Aegypti coluerunt deos alienos 17:8 et am- 17:8 And walked in the statutes of the heathen, whom the

LORD cast out from before the children of Israel, and of the kings of Israel, which they had made.

bulauerunt iuxta ritum gentium quas consumpserat Dominus in conspectu filiorum Israhel et regum Israhel quia similiter fecerant 17:9 et operuerunt filii Israhel uerbis non re- 17:9 And the children of Israel did secretly those things

that were not right against the LORD their God, and they built them high places in all their cities, from the tower of

ctis Dominum Deum suum et aedificauerunt sibi excelsa in the watchmen to the fenced city. cunctis urbibus suis a turre custodum usque ad ciuitatem munitam 17:10 feceruntque sibi statuas et lucos in omni colle 17:10 And they set them up images and groves in every

high hill, and under every green tree:

sublimi et subter omne lignum nemorosum 17:11 et adole- 17:11 And there they burnt incense in all the high places, as did the heathen whom the LORD carried away before them; and wrought wicked things to provoke the LORD to

bant ibi incensum super aras in more gentium quas trans- anger: tulerat Dominus a facie eorum feceruntque uerba pessima inritantes Dominum 17:12 et coluerunt inmunditias de qui- 17:12 For they served idols, whereof the LORD had said

unto them, Ye shall not do this thing.

bus praecepit Dominus eis ne facerent uerbum hoc 17:13 17:13 Yet the LORD testified against Israel, and against Judah, by all the prophets, and by all the seers, saying, Turn ye from your evil ways, and keep my commandments and my statutes, according to all the law which I commanded your fathers, and which I sent to you by my servants the prophets.

et testificatus est Dominus in Israhel et in Iuda per manum

Biblia Sacra Vulgata

omnium prophetarum et uidentum dicens reuertimini a uiis uestris pessimis et custodite praecepta mea et caerimonias iuxta omnem legem quam praecepi patribus uestris et sicut

17:14 Notwithstanding they would not hear, but hardened misi ad uos in manu seruorum meorum prophetarum 17:14 their necks, like to the neck of their fathers, that did not believe in the LORD their God.

qui non audierunt sed indurauerunt ceruicem suam iuxta ceruicem patrum suorum qui noluerunt oboedire Domino Deo

17:15 And they rejected his statutes, and his covenant that suo 17:15 et abiecerunt legitima eius et pactum quod pepihe made with their fathers, and his testimonies which he testified against them; and they followed vanity, and became vain, and went after the heathen that were round about them, concerning whom the LORD had charged them, that they should not do like them.

git cum patribus eorum et testificationes quibus contestatus est eos secutique sunt uanitates et uane egerunt et secuti sunt gentes quae erant per circuitum eorum super quibus praeceperat Dominus eis ut non facerent sicut et illae faciebant

17:16 And they left all the commandments of the LORD 17:16 et dereliquerunt omnia praecepta Domini Dei sui fe-

their God, and made them molten images, even two calves, and made a grove, and worshipped all the host of heaven,

and served Baal. ceruntque sibi conflatiles duos uitulos et lucos et adoraue-17:17 And they caused their sons and their daughters to runt uniuersam militiam caeli seruieruntque Baal 17:17 et

pass through the fire, and used divination and enchantments, and sold themselves to do evil in the sight of the

LORD, to provoke him to anger. consecrabant ei filios suos et filias suas per ignem et diuina-

tionibus inseruiebant et auguriis et tradiderunt se ut facerent

17:18 Therefore the LORD was very angry with Israel, malum coram Domino et inritarent eum 17:18 iratusque est

and removed them out of his sight: there was none left but the tribe of Judah only.

Dominus uehementer Israhel et abstulit eos de conspectu suo

17:19 Also Judah kept not the commandments of the et non remansit nisi tribus Iuda tantummodo 17:19 sed nec

LORD their God, but walked in the statutes of Israel which they made.

ipse Iuda custodiuit mandata Domini Dei sui uerum ambu-

17:20 And the LORD rejected all the seed of Israel, and lauit in erroribus Israhel quos operatus fuerat 17:20 proie-

afflicted them, and delivered them into the hand of spoilers, until he had cast them out of his sight.

citque Dominus omne semen Israhel et adflixit eos et tradidit in manu diripientium donec proiceret eos a facie sua

17:21 For he rent Israel from the house of David; and 17:21 ex eo iam tempore quo scissus est Israhel a domo Da-

they made Jeroboam the son of Nebat king: and Jeroboam drave Israel from following the LORD, and made them sin

a great sin. uid et constituerunt sibi regem Hieroboam filium Nabath se-

parauit enim Hieroboam Israhel a Domino et peccare eos fe-

17:22 For the children of Israel walked in all the sins of cit peccatum magnum 17:22 et ambulauerunt filii Israhel in

Jeroboam which he did; they departed not from them;

uniuersis peccatis Hieroboam quae fecerat non recesserunt

17:23 Until the LORD removed Israel out of his sight, as ab eis 17:23 usquequo auferret Dominus Israhel a facie sua

he had said by all his servants the prophets. So was Israel carried away out of their own land to Assyria unto this day.

sicut locutus fuerat in manu omnium seruorum suorum prophetarum translatusque est Israhel de terra sua in Assyrios

17:24 And the king of Assyria brought men from Baby- usque in diem hanc 17:24 adduxit autem rex Assyriorum

lon, and from Cuthah, and from Ava, and from Hamath, and from Sepharvaim, and placed them in the cities of Samaria instead of the children of Israel: and they possessed Samaria, and dwelt in the cities thereof.

de Babylone et de Chutha et de Haiath et de Emath et de Se-

Liber II Regum

pharuaim et conlocauit eos in ciuitatibus Samariae pro filiis Israhel qui possederunt Samariam et habitauerunt in urbibus eius 17:25 cumque ibi habitare coepissent non timebant 17:25 And so it was at the beginning of their dwelling

there, that they feared not the LORD: therefore the LORD sent lions among them, which slew some of them.

Dominum et inmisit eis Dominus leones qui interficiebant eos 17:26 nuntiatumque est regi Assyriorum et dictum gen- 17:26 Wherefore they spake to the king of Assyria, say-

ing, The nations which thou hast removed, and placed in the cities of Samaria, know not the manner of the God of the land: therefore he hath sent lions among them, and, behold, they slay them, because they know not the manner of the God of the land.

tes quas transtulisti et habitare fecisti in ciuitatibus Samariae ignorant legitima Dei terrae et inmisit in eos Dominus leones et ecce interficiunt eos eo quod ignorent ritum Dei terrae 17:27 praecepit autem rex Assyriorum dicens ducite 17:27 Then the king of Assyria commanded, saying, Carry

thither one of the priests whom ye brought from thence; and let them go and dwell there, and let him teach them

illuc unum de sacerdotibus quos inde captiuos adduxistis et the manner of the God of the land. uadat et habitet cum eis et doceat eos legitima Dei terrae 17:28 igitur cum uenisset unus de sacerdotibus his qui cap- 17:28 Then one of the priests whom they had carried away

from Samaria came and dwelt in Bethel, and taught them how they should fear the LORD.

tiui ducti fuerant de Samaria habitauit in Bethel et docebat eos quomodo colerent Dominum 17:29 et unaquaeque gens 17:29 Howbeit every nation made gods of their own, and

put them in the houses of the high places which the Samaritans had made, every nation in their cities wherein they

fabricata est deum suum posueruntque eos in fanis excel- dwelt. sis quae fecerant Samaritae gens et gens in urbibus suis in quibus habitabant 17:30 uiri enim babylonii fecerunt Soc- 17:30 And the men of Babylon made Succothbenoth, and

the men of Cuth made Nergal, and the men of Hamath made Ashima,

chothbenoth uiri autem chutheni fecerunt Nergel et uiri de Emath fecerunt Asima 17:31 porro Euei fecerunt Nebaaz 17:31 And the Avites made Nibhaz and Tartak, and the

Sepharvites burnt their children in fire to Adrammelech and Anammelech, the gods of Sepharvaim.

et Tharthac hii autem qui erant de Sepharuaim conburebant filios suos igni Adramelech et Anamelech diis Sepharuaim 17:32 et nihilominus colebant Dominum fecerunt autem sibi 17:32 So they feared the LORD, and made unto them-

selves of the lowest of them priests of the high places, which sacrificed for them in the houses of the high places.

de nouissimis sacerdotes excelsorum et ponebant eos in fanis sublimibus 17:33 et cum Dominum colerent diis quo- 17:33 They feared the LORD, and served their own gods,

after the manner of the nations whom they carried away from thence.

que suis seruiebant iuxta consuetudinem gentium de quibus translati fuerant Samariam 17:34 usque in praesentem diem 17:34 Unto this day they do after the former manners: they

fear not the LORD, neither do they after their statutes, or after their ordinances, or after the law and commandment which the LORD commanded the children of Jacob, whom he named Israel;

morem sequuntur antiquum non timent Dominum neque custodiunt caerimonias eius et iudicia et legem et mandatum quod praeceperat Dominus filiis Iacob quem cognominauit Israhel 17:35 et percusserat cum eis pactum et mandaue- 17:35 With whom the LORD had made a covenant, and

charged them, saying, Ye shall not fear other gods, nor bow yourselves to them, nor serve them, nor sacrifice to

rat eis dicens nolite timere deos alienos et non adoretis eos them: neque colatis et non immoletis eis 17:36 sed Dominum De- 17:36 But the LORD, who brought you up out of the land

of Egypt with great power and a stretched out arm, him shall ye fear, and him shall ye worship, and to him shall

um uestrum qui eduxit uos de terra Aegypti in fortitudine ye do sacrifice.

Biblia Sacra Vulgata

magna et in brachio extento ipsum timete illum adorate et

17:37 And the statutes, and the ordinances, and the law, ipsi immolate 17:37 caerimonias quoque et iudicia et legem

and the commandment, which he wrote for you, ye shall observe to do for evermore; and ye shall not fear other

gods. et mandatum quod scripsit uobis custodite ut faciatis cunctis 17:38 And the covenant that I have made with you ye shall diebus et non timeatis deos alienos 17:38 et pactum quod

not forget; neither shall ye fear other gods.

percussi uobiscum nolite obliuisci nec colatis deos alienos

17:39 But the LORD your God ye shall fear; and he shall 17:39 sed Dominum Deum uestrum timete et ipse eruet uos

deliver you out of the hand of all your enemies.

17:40 Howbeit they did not hearken, but they did after de manu omnium inimicorum uestrorum 17:40 illi uero non

their former manner.

audierunt sed iuxta consuetudinem suam pristinam perpetra-

17:41 So these nations feared the LORD, and served their bant 17:41 fuerunt igitur gentes istae timentes quidem Do-

graven images, both their children, and their children’s children: as did their fathers, so do they unto this day.

minum sed nihilominus et idolis suis seruientes nam et filii eorum et nepotes sicut fecerunt parentes sui ita faciunt usque in praesentem diem

18:1 Now it came to pass in the third year of Hoshea son 18:1 anno tertio Osee filii Hela regis Israhel regnauit Eze-

of Elah king of Israel, that Hezekiah the son of Ahaz king of Judah began to reign.

chias filius Ahaz regis Iuda 18:2 uiginti quinque annorum

18:2 Twenty and five years old was he when he began to reign; and he reigned twenty and nine years in Jerusalem. His mother’s name also was Abi, the daughter of Zachariah.

erat cum regnare coepisset et uiginti et nouem annis regnauit in Hierusalem nomen matris eius Abi filia Zacchariae

18:3 And he did that which was right in the sight of the 18:3 fecitque quod erat bonum coram Domino iuxta omnia

LORD, according to all that David his father did.

18:4 He removed the high places, and brake the images, quae fecerat Dauid pater suus 18:4 ipse dissipauit excelsa

and cut down the groves, and brake in pieces the brasen serpent that Moses had made: for unto those days the children of Israel did burn incense to it: and he called it Nehushtan.

et contriuit statuas et succidit lucos confregitque serpentem aeneum quem fecerat Moses siquidem usque ad illud tempus filii Israhel adolebant ei incensum uocauitque eum Naasthan

18:5 He trusted in the LORD God of Israel; so that after 18:5 in Domino Deo Israhel sperauit itaque post eum non fuit

him was none like him among all the kings of Judah, nor any that were before him.

similis ei de cunctis regibus Iuda sed neque in his qui ante

18:6 For he clave to the LORD, and departed not from eum fuerunt 18:6 et adhesit Domino et non recessit a ues-

following him, but kept his commandments, which the LORD commanded Moses.

tigiis eius fecitque mandata eius quae praeceperat Dominus

18:7 And the LORD was with him; and he prospered Mosi 18:7 unde et erat Dominus cum eo et in cunctis ad

whithersoever he went forth: and he rebelled against the king of Assyria, and served him not.

quae procedebat sapienter se agebat rebellauit quoque con-

18:8 He smote the Philistines, even unto Gaza, and the tra regem Assyriorum et non seruiuit ei 18:8 ipse percussit

borders thereof, from the tower of the watchmen to the fenced city.

Philistheos usque Gazam et omnes terminos eorum a turre

18:9 And it came to pass in the fourth year of king custodum usque ad ciuitatem muratam 18:9 anno quarto

Hezekiah, which was the seventh year of Hoshea son of Elah king of Israel, that Shalmaneser king of Assyria came

up against Samaria, and besieged it. regis Ezechiae qui erat annus septimus Osee filii Hela re-

gis Israhel ascendit Salmanassar rex Assyriorum Samariam

18:10 And at the end of three years they took it: even in the et obpugnauit eam 18:10 et cepit nam post annos tres anno sixth year of Hezekiah, that is in the ninth year of Hoshea king of Israel, Samaria was taken.

Liber II Regum

sexto Ezechiae id est nono anno Osee regis Israhel capta est Samaria 18:11 et transtulit rex Assyriorum Israhel in As- 18:11 And the king of Assyria did carry away Israel unto

Assyria, and put them in Halah and in Habor by the river of Gozan, and in the cities of the Medes:

syrios conlocauitque eos in Ala et in Habor fluuiis Gozan in ciuitatibus Medorum 18:12 quia non audierunt uocem 18:12 Because they obeyed not the voice of the LORD

their God, but transgressed his covenant, and all that Moses the servant of the LORD commanded, and would

Domini Dei sui sed praetergressi sunt pactum eius omnia not hear them, nor do them. quae praeceperat Moses seruus Domini non audierunt neque fecerunt 18:13 anno quartodecimo regis Ezechiae ascendit 18:13 Now in the fourteenth year of king Hezekiah did

Sennacherib king of Assyria come up against all the fenced cities of Judah, and took them.

Sennacherib rex Assyriorum ad uniuersas ciuitates Iuda munitas et cepit eas 18:14 tunc misit Ezechias rex Iuda nun- 18:14 And Hezekiah king of Judah sent to the king of As-

syria to Lachish, saying, I have offended; return from me: that which thou puttest on me will I bear. And the king of Assyria appointed unto Hezekiah king of Judah three hundred talents of silver and thirty talents of gold.

tios ad regem Assyriorum Lachis dicens peccaui recede a me et omne quod inposueris mihi feram indixit itaque rex Assyriorum Ezechiae regi Iudae trecenta talenta argenti et triginta talenta auri 18:15 deditque Ezechias omne argen- 18:15 And Hezekiah gave him all the silver that was found

in the house of the LORD, and in the treasures of the king’s house.

tum quod reppertum fuerat in domo Domini et in thesauris regis 18:16 in tempore illo confregit Ezechias ualuas templi 18:16 At that time did Hezekiah cut off the gold from the

doors of the temple of the LORD, and from the pillars which Hezekiah king of Judah had overlaid, and gave it to

Domini et lamminas auri quas ipse adfixerat et dedit eas re- the king of Assyria. gi Assyriorum 18:17 misit autem rex Assyriorum Tharthan 18:17 And the king of Assyria sent Tartan and Rabsaris

and Rabshakeh from Lachish to king Hezekiah with a great host against Jerusalem. And they went up and came to Jerusalem. And when they were come up, they came and stood by the conduit of the upper pool, which is in the highway of the fuller’s field.

et Rabsaris et Rabsacen de Lachis ad regem Ezechiam cum manu ualida Hierusalem qui cum ascendissent uenerunt in Hierusalem et steterunt iuxta aquaeductum piscinae superioris quae est in uia agri Fullonis 18:18 uocaueruntque regem 18:18 And when they had called to the king, there came

out to them Eliakim the son of Hilkiah, which was over the household, and Shebna the scribe, and Joah the son of

egressus est autem ad eos Eliachim filius Helciae praepositus Asaph the recorder. domus et Sobna scriba et Ioahe filius Asaph a commentariis 18:19 dixitque ad eos Rabsaces loquimini Ezechiae haec dicit 18:19 And Rabshakeh said unto them, Speak ye now to

Hezekiah, Thus saith the great king, the king of Assyria, What confidence is this wherein thou trustest?

rex magnus rex Assyriorum quae est ista fiducia qua niteris 18:20 forsitan inisti consilium ut praepares te ad proelium in 18:20 Thou sayest, (but they are but vain words,) I have

counsel and strength for the war. Now on whom dost thou trust, that thou rebellest against me?

quo confidis ut audeas rebellare 18:21 an speras in baculo

18:21 Now, behold, thou trustest upon the staff of this bruised reed, even upon Egypt, on which if a man lean, it will go into his hand, and pierce it: so is Pharaoh king of Egypt unto all that trust on him.

harundineo atque confracto Aegypto super quem si incubuerit homo comminutus ingreditur manum eius et perforabit eam sic est Pharao rex Aegypti omnibus qui confidunt in se 18:22 quod si dixeritis mihi in Domino Deo nostro habemus 18:22 But if ye say unto me, We trust in the LORD our

God: is not that he, whose high places and whose altars Hezekiah hath taken away, and hath said to Judah and

fiduciam nonne iste est cuius abstulit Ezechias excelsa et al- Jerusalem, Ye shall worship before this altar in Jerusalem? taria et praecepit Iudae et Hierusalem ante altare hoc adora-

Biblia Sacra Vulgata

18:23 Now therefore, I pray thee, give pledges to my lord bitis in Hierusalem 18:23 nunc igitur transite ad dominum

the king of Assyria, and I will deliver thee two thousand horses, if thou be able on thy part to set riders upon them.

meum regem Assyriorum et dabo uobis duo milia equorum

18:24 How then wilt thou turn away the face of one captain et uidete an habere ualeatis ascensores eorum 18:24 et quo-

of the least of my master’s servants, and put thy trust on Egypt for chariots and for horsemen?

modo potestis resistere ante unum satrapam de seruis domini mei minimis an fiduciam habes in Aegypto propter currus et

18:25 Am I now come up without the LORD against this equites 18:25 numquid sine Domini uoluntate ascendi ad

place to destroy it? The LORD said to me, Go up against this land, and destroy it.

locum istum ut demolirer eum Dominus dixit mihi ascende

18:26 Then said Eliakim the son of Hilkiah, and Shebna, ad terram hanc et demolire eam 18:26 dixerunt autem Elia-

and Joah, unto Rabshakeh, Speak, I pray thee, to thy servants in the Syrian language; for we understand it: and talk not with us in the Jews’ language in the ears of the people that are on the wall.

chim filius Helciae et Sobna et Ioahe Rabsaci precamur ut loquaris nobis seruis tuis syriace siquidem intellegimus hanc linguam et non loquaris nobis iudaice audiente populo qui

18:27 But Rabshakeh said unto them, Hath my master sent est super murum 18:27 responditque eis Rabsaces numquid

me to thy master, and to thee, to speak these words? hath he not sent me to the men which sit on the wall, that they may eat their own dung, and drink their own piss with you?

ad dominum tuum et ad te misit me dominus meus ut loquerer sermones hos et non ad uiros qui sedent super murum ut comedant stercora sua et bibant urinam suam uobiscum

18:28 Then Rabshakeh stood and cried with a loud voice 18:28 stetit itaque Rabsaces et clamauit uoce magna iudaice

in the Jews’ language, and spake, saying, Hear the word of the great king, the king of Assyria:

et ait audite uerba regis magni regis Assyriorum 18:29 haec

18:29 Thus saith the king, Let not Hezekiah deceive you: for he shall not be able to deliver you out of his hand:

dicit rex non uos seducat Ezechias non enim poterit eruere

18:30 Neither let Hezekiah make you trust in the LORD, uos de manu mea 18:30 neque fiduciam uobis tribuat susaying, The LORD will surely deliver us, and this city shall not be delivered into the hand of the king of Assyria.

per Domino dicens eruens liberabit nos Dominus et non tra-

18:31 Hearken not to Hezekiah: for thus saith the king of detur ciuitas haec in manu regis Assyriorum 18:31 nolite

Assyria, Make an agreement with me by a present, and come out to me, and then eat ye every man of his own vine, and every one of his fig tree, and drink ye every one the waters of his cistern:

audire Ezechiam haec enim dicit rex Assyriorum facite mecum quod uobis est utile et egredimini ad me et comedet unusquisque de uinea sua et de ficu sua et bibetis aquas de

18:32 Until I come and take you away to a land like your cisternis uestris 18:32 donec ueniam et transferam uos in

own land, a land of corn and wine, a land of bread and vineyards, a land of oil olive and of honey, that ye may live, and not die: and hearken not unto Hezekiah, when he persuadeth you, saying, The LORD will deliver us.

terram quae similis terrae uestrae est in terram fructiferam et fertilem uini terram panis et uinearum terram oliuarum et olei ac mellis et uiuetis et non moriemini nolite audire Eze-

18:33 Hath any of the gods of the nations delivered at all chiam qui uos decipit dicens Dominus liberabit nos 18:33

his land out of the hand of the king of Assyria?

numquid liberauerunt dii gentium terram suam de manu re-

18:34 Where are the gods of Hamath, and of Arpad? gis Assyriorum 18:34 ubi est deus Emath et Arfad ubi est

where are the gods of Sepharvaim, Hena, and Ivah? have they delivered Samaria out of mine hand?

deus Sepharuaim Ana et Aua numquid liberauerunt Sama-

18:35 Who are they among all the gods of the countries, riam de manu mea 18:35 quinam illi sunt in uniuersis diis

that have delivered their country out of mine hand, that the LORD should deliver Jerusalem out of mine hand?

Liber II Regum

terrarum qui eruerunt regionem suam de manu mea ut possit eruere Dominus Hierusalem de manu mea 18:36 tacuit ita- 18:36 But the people held their peace, and answered him

not a word: for the king’s commandment was, saying, Answer him not.

que populus et non respondit ei quicquam siquidem praeceptum regis acceperant ut non responderent ei 18:37 uenitque 18:37 Then came Eliakim the son of Hilkiah, which was

over the household, and Shebna the scribe, and Joah the son of Asaph the recorder, to Hezekiah with their clothes

Eliachim filius Helciae praepositus domus et Sobna scriba et rent, and told him the words of Rabshakeh. Ioahe filius Asaph a commentariis ad Ezechiam scissis uestibus et nuntiauerunt ei uerba Rabsacis

19:1 quae cum audisset rex Ezechias scidit uestimenta sua 19:1 And it came to pass, when king Hezekiah heard it,

that he rent his clothes, and covered himself with sackcloth, and went into the house of the LORD.

et opertus est sacco ingressusque est domum Domini 19:2

19:2 And he sent Eliakim, which was over the household, and Shebna the scribe, and the elders of the priests, covered with sackcloth, to Isaiah the prophet the son of Amoz.

et misit Eliachim praepositum domus et Sobnam scribam et senes de sacerdotibus opertos saccis ad Esaiam prophetam filium Amos 19:3 qui dixerunt haec dicit Ezechias dies 19:3 And they said unto him, Thus saith Hezekiah, This

day is a day of trouble, and of rebuke, and blasphemy; for the children are come to the birth, and there is not strength

tribulationis et increpationis et blasphemiae dies iste uene- to bring forth. runt filii usque ad partum et uires non habet parturiens 19:4 19:4 It may be the LORD thy God will hear all the words

of Rabshakeh, whom the king of Assyria his master hath sent to reproach the living God; and will reprove the words which the LORD thy God hath heard: wherefore lift up thy prayer for the remnant that are left.

si forte audiat Dominus Deus tuus uniuersa uerba Rabsacis quem misit rex Assyriorum dominus suus ut exprobraret Deum uiuentem et argueret uerbis quae audiuit Dominus Deus tuus et fac orationem pro reliquiis quae reppertae sunt 19:5 uenerunt ergo serui regis Ezechiae ad Esaiam 19:5 So the servants of king Hezekiah came to Isaiah. 19:6 dixitque eis Esaias haec dicetis domino uestro haec dicit 19:6 And Isaiah said unto them, Thus shall ye say to your

master, Thus saith the LORD, Be not afraid of the words which thou hast heard, with which the servants of the king

Dominus noli timere a facie sermonum quos audisti quibus of Assyria have blasphemed me. blasphemauerunt pueri regis Assyriorum me 19:7 ecce ego 19:7 Behold, I will send a blast upon him, and he shall

hear a rumour, and shall return to his own land; and I will cause him to fall by the sword in his own land.

inmittam ei spiritum et audiet nuntium et reuertetur in terram suam et deiciam eum gladio in terra sua 19:8 reuersus 19:8 So Rabshakeh returned, and found the king of As-

syria warring against Libnah: for he had heard that he was departed from Lachish.

est igitur Rabsaces et inuenit regem Assyriorum expugnantem Lobnam audierat enim quod recessisset de Lachis 19:9 19:9 And when he heard say of Tirhakah king of Ethiopia,

Behold, he is come out to fight against thee: he sent messengers again unto Hezekiah, saying,

cumque audisset de Tharaca rege Aethiopiae dicentes ecce egressus est ut pugnet aduersum te et iret contra eum misit nuntios ad Ezechiam dicens 19:10 haec dicite Ezechiae 19:10 Thus shall ye speak to Hezekiah king of Judah, say-

ing, Let not thy God in whom thou trustest deceive thee, saying, Jerusalem shall not be delivered into the hand of

regi Iudae non te seducat Deus tuus in quo habes fiduciam the king of Assyria. neque dicas non tradetur Hierusalem in manu regis Assyriorum 19:11 tu enim ipse audisti quae fecerint reges Assyrio- 19:11 Behold, thou hast heard what the kings of Assyria

have done to all lands, by destroying them utterly: and shalt thou be delivered?

rum uniuersis terris quomodo uastauerint eas num ergo solus

Biblia Sacra Vulgata

19:12 Have the gods of the nations delivered them which poteris liberari 19:12 numquid liberauerunt dii gentium sin-

my fathers have destroyed; as Gozan, and Haran, and Rezeph, and the children of Eden which were in Thelasar?

gulos quos uastauerunt patres mei Gozan uidelicet et Aran

19:13 Where is the king of Hamath, and the king of Arpad, et Reseph et filios Eden qui erant in Thelassar 19:13 ubi est

and the king of the city of Sepharvaim, of Hena, and Ivah?

rex Emath et rex Arfad et rex ciuitatis Sepharuaim Ana et

19:14 And Hezekiah received the letter of the hand of the Aua 19:14 itaque cum accepisset Ezechias litteras de manu

messengers, and read it: and Hezekiah went up into the house of the LORD, and spread it before the LORD.

nuntiorum et legisset eas ascendit in domum Domini et ex-

19:15 And Hezekiah prayed before the LORD, and said, O pandit eas coram Domino 19:15 et orauit in conspectu eius

LORD God of Israel, which dwellest between the cherubims, thou art the God, even thou alone, of all the king-

doms of the earth; thou hast made heaven and earth. dicens Domine Deus Israhel qui sedes super cherubin tu es

Deus solus regum omnium terrae tu fecisti caelum et ter-

19:16 LORD, bow down thine ear, and hear: open, LORD, ram 19:16 inclina aurem tuam et audi aperi Domine oculos

thine eyes, and see: and hear the words of Sennacherib, which hath sent him to reproach the living God.

tuos et uide et audi omnia uerba Sennacherib qui misit ut

19:17 Of a truth, LORD, the kings of Assyria have de- exprobraret nobis Deum uiuentem 19:17 uere Domine dis-

stroyed the nations and their lands,

19:18 And have cast their gods into the fire: for they were sipauerunt reges Assyriorum gentes et terras omnium 19:18

no gods, but the work of men’s hands, wood and stone: therefore they have destroyed them.

et miserunt deos eorum in ignem non enim erant dii sed opera manuum hominum e ligno et lapide et perdiderunt

19:19 Now therefore, O LORD our God, I beseech thee, eos 19:19 nunc igitur Domine Deus noster saluos nos fac

save thou us out of his hand, that all the kingdoms of the earth may know that thou art the LORD God, even thou

only. de manu eius ut sciant omnia regna terrae quia tu es Do-

19:20 Then Isaiah the son of Amoz sent to Hezekiah, say- minus Deus solus 19:20 misit autem Esaias filius Amos ad ing, Thus saith the LORD God of Israel, That which thou hast prayed to me against Sennacherib king of Assyria I

have heard. Ezechiam dicens haec dicit Dominus Deus Israhel quae de-

precatus es me super Sennacherib rege Assyriorum audiui

19:21 This is the word that the LORD hath spoken con- 19:21 iste est sermo quem locutus est Dominus de eo spreuit

cerning him; The virgin the daughter of Zion hath despised thee, and laughed thee to scorn; the daughter of Jerusalem

hath shaken her head at thee. te et subsannauit uirgo filia Sion post tergum tuum caput mo-19:22 Whom hast thou reproached and blasphemed? and uit filia Hierusalem 19:22 cui exprobrasti et quem blasphe-

against whom hast thou exalted thy voice, and lifted up thine eyes on high? even against the Holy One of Israel.

masti contra quem exaltasti uocem et eleuasti in excelsum

19:23 By thy messengers thou hast reproached the LORD, oculos tuos contra Sanctum Israhel 19:23 per manum ser-

and hast said, With the multitude of my chariots I am come up to the height of the mountains, to the sides of Lebanon, and will cut down the tall cedar trees thereof, and the choice fir trees thereof: and I will enter into the lodgings of his borders, and into the forest of his Carmel.

uorum tuorum exprobrasti Domino et dixisti in multitudine curruum meorum ascendi excelsa montium in summitate Libani et succidi sublimes cedros eius electas abietes eius et ingressus sum usque ad terminos eius saltum Carmeli eius

19:24 I have digged and drunk strange waters, and with the 19:24 ego succidi et bibi aquas alienas et siccaui uestigiis

sole of my feet have I dried up all the rivers of besieged places.

pedum meorum omnes aquas clausas 19:25 numquid non

19:25 Hast thou not heard long ago how I have done it, and of ancient times that I have formed it? now have I brought it to pass, that thou shouldest be to lay waste fenced cities into ruinous heaps.

audisti quid ab initio fecerim ex diebus antiquis plasmaui illud et nunc adduxi eruntque in ruinam collium pugnantium

Liber II Regum

ciuitates munitae 19:26 et qui sedent in eis humiles manu 19:26 Therefore their inhabitants were of small power,

they were dismayed and confounded; they were as the grass of the field, and as the green herb, as the grass on

contremuerunt et confusi sunt facti sunt quasi faenum agri et the house tops, and as corn blasted before it be grown up. uirens herba tectorum quae arefacta est antequam ueniret ad maturitatem 19:27 habitaculum tuum et egressum tuum et 19:27 But I know thy abode, and thy going out, and thy

coming in, and thy rage against me.

uiam tuam ego praesciui et furorem tuum contra me 19:28 19:28 Because thy rage against me and thy tumult is come

up into mine ears, therefore I will put my hook in thy nose, and my bridle in thy lips, and I will turn thee back by the

insanisti in me et superbia tua ascendit in aures meas po- way by which thou camest. nam itaque circulum in naribus tuis et camum in labris tuis

et reducam te in uiam per quam uenisti 19:29 tibi autem 19:29 And this shall be a sign unto thee, Ye shall eat this year such things as grow of themselves, and in the second year that which springeth of the same; and in the third year sow ye, and reap, and plant vineyards, and eat the fruits thereof.

Ezechia hoc erit signum comede hoc anno quod reppereris in secundo autem anno quae sponte nascuntur porro in anno tertio seminate et metite plantate uineas et comedite fructum earum 19:30 et quodcumque reliquum fuerit de domo Iuda 19:30 And the remnant that is escaped of the house of Ju-

dah shall yet again take root downward, and bear fruit upward.

mittet radicem deorsum et faciet fructum sursum 19:31 de

19:31 For out of Jerusalem shall go forth a remnant, and they that escape out of mount Zion: the zeal of the LORD of hosts shall do this.

Hierusalem quippe egredientur reliquiae et quod saluetur

de monte Sion zelus Domini exercituum faciet hoc 19:32 19:32 Therefore thus saith the LORD concerning the king of Assyria, He shall not come into this city, nor shoot an arrow there, nor come before it with shield, nor cast a bank

quam ob rem haec dicit Dominus de rege Assyriorum non against it. ingredietur urbem hanc nec mittet in eam sagittam nec occupabit eam clypeus nec circumdabit eam munitio 19:33 per 19:33 By the way that he came, by the same shall he re-

turn, and shall not come into this city, saith the LORD.

uiam qua uenit reuertetur et ciuitatem hanc non ingredietur

dicit Dominus 19:34 protegamque urbem hanc et saluabo 19:34 For I will defend this city, to save it, for mine own

sake, and for my servant David’s sake.

eam propter me et propter Dauid seruum meum 19:35 fac- 19:35 And it came to pass that night, that the angel of

the LORD went out, and smote in the camp of the Assyrians an hundred fourscore and five thousand: and when they arose early in the morning, behold, they were all dead corpses.

tum est igitur in nocte illa uenit angelus Domini et percussit castra Assyriorum centum octoginta quinque milia cumque diluculo surrexisset uidit omnia corpora mortuorum et recedens abiit 19:36 et reuersus est Sennacherib rex Assyrio- 19:36 So Sennacherib king of Assyria departed, and went

and returned, and dwelt at Nineveh.

rum et mansit in Nineue 19:37 cumque adoraret in templo 19:37 And it came to pass, as he was worshipping in the

house of Nisroch his god, that Adrammelech and Sharezer his sons smote him with the sword: and they escaped into the land of Armenia. And Esarhaddon his son reigned in his stead.

Neserach deum suum Adramelech et Sarasar filii eius percusserunt eum gladio fugeruntque in terram Armeniorum et regnauit Eseraddon filius eius pro eo

20:1 in diebus illis aegrotauit Ezechias usque ad mortem 20:1 In those days was Hezekiah sick unto death. And the

prophet Isaiah the son of Amoz came to him, and said unto him, Thus saith the LORD, Set thine house in order; for

et uenit ad eum Esaias filius Amos prophetes dixitque ei thou shalt die, and not live. haec dicit Dominus Deus praecipe domui tuae morieris enim et non uiues 20:2 qui conuertit faciem suam ad parietem 20:2 Then he turned his face to the wall, and prayed unto

the LORD, saying,

Biblia Sacra Vulgata

20:3 I beseech thee, O LORD, remember now how I have et orauit Dominum dicens 20:3 obsecro Domine memento

walked before thee in truth and with a perfect heart, and have done that which is good in thy sight. And Hezekiah

wept sore. quomodo ambulauerim coram te in ueritate et in corde per-

fecto et quod placitum est coram te fecerim fleuit itaque Eze-

20:4 And it came to pass, afore Isaiah was gone out into chias fletu magno 20:4 et antequam egrederetur Esaias me-

the middle court, that the word of the LORD came to him, saying,

diam partem atrii factus est sermo Domini ad eum dicens

20:5 Turn again, and tell Hezekiah the captain of my peo- 20:5 reuertere et dic Ezechiae duci populi mei haec dicit Do-

ple, Thus saith the LORD, the God of David thy father, I have heard thy prayer, I have seen thy tears: behold, I will heal thee: on the third day thou shalt go up unto the house of the LORD.

minus Deus Dauid patris tui audiui orationem tuam uidi lacrimam tuam et ecce sanaui te die tertio ascendes templum

20:6 And I will add unto thy days fifteen years; and I will Domini 20:6 et addam diebus tuis quindecim annos sed et deliver thee and this city out of the hand of the king of Assyria; and I will defend this city for mine own sake, and

for my servant David’s sake. de manu regis Assyriorum liberabo te et ciuitatem hanc et

protegam urbem istam propter me et propter Dauid seruum

20:7 And Isaiah said, Take a lump of figs. And they took meum 20:7 dixitque Esaias adferte massam ficorum quam

and laid it on the boil, and he recovered.

cum adtulissent et posuissent super ulcus eius curatus est

20:8 And Hezekiah said unto Isaiah, What shall be the sign 20:8 dixerat autem Ezechias ad Esaiam quod erit signum quia

that the LORD will heal me, and that I shall go up into the house of the LORD the third day?

Dominus me sanabit et quia ascensurus sum die tertio tem-

20:9 And Isaiah said, This sign shalt thou have of the plum Domini 20:9 cui ait Esaias hoc erit signum a Domino

LORD, that the LORD will do the thing that he hath spoken: shall the shadow go forward ten degrees, or go back

ten degrees? quod facturus sit Dominus sermonem quem locutus est uis ut

accedat umbra decem lineis an ut reuertatur totidem gradibus

20:10 And Hezekiah answered, It is a light thing for the 20:10 et ait Ezechias facile est umbram crescere decem lineis

shadow to go down ten degrees: nay, but let the shadow return backward ten degrees.

nec hoc uolo ut fiat sed ut reuertatur retrorsum decem gra-

20:11 And Isaiah the prophet cried unto the LORD: and dibus 20:11 inuocauit itaque Esaias propheta Dominum et

he brought the shadow ten degrees backward, by which it had gone down in the dial of Ahaz.

reduxit umbram per lineas quibus iam descenderat in horo-

20:12 At that time Berodachbaladan, the son of Baladan, logio Ahaz retrorsum decem gradibus 20:12 in tempore illo

king of Babylon, sent letters and a present unto Hezekiah: for he had heard that Hezekiah had been sick.

misit Berodach Baladan filius Baladan rex Babyloniorum litteras et munera ad Ezechiam audierat enim quod aegrotasset

20:13 And Hezekiah hearkened unto them, and shewed Ezechias 20:13 laetatus est autem in aduentum eorum Eze-

them all the house of his precious things, the silver, and the gold, and the spices, and the precious ointment, and all the house of his armour, and all that was found in his treasures: there was nothing in his house, nor in all his dominion, that Hezekiah shewed them not.

chias et ostendit eis domum aromatum et aurum et argentum et pigmenta uaria unguenta quoque et domum uasorum suorum et omnia quae habere potuerat in thesauris suis non fuit quod non monstraret eis Ezechias in domo sua et in omni

20:14 Then came Isaiah the prophet unto king Hezekiah, potestate sua 20:14 uenit autem Esaias propheta ad regem

and said unto him, What said these men? and from whence came they unto thee? And Hezekiah said, They are come

from a far country, even from Babylon. Ezechiam dixitque ei quid dixerunt uiri isti aut unde uenerunt ad te cui ait Ezechias de terra longinqua uenerunt de

Liber II Regum

Babylone 20:15 at ille respondit quid uiderunt in domo tua 20:15 And he said, What have they seen in thine house?

And Hezekiah answered, All the things that are in mine house have they seen: there is nothing among my treasures

ait Ezechias omnia quae sunt in domo mea uiderunt nihil that I have not shewed them. est quod non monstrauerim eis in thesauris meis 20:16 dixit 20:16 And Isaiah said unto Hezekiah, Hear the word of

the LORD.

itaque Esaias Ezechiae audi sermonem Domini 20:17 ecce 20:17 Behold, the days come, that all that is in thine house, and that which thy fathers have laid up in store unto this day, shall be carried into Babylon: nothing shall be left,

dies uenient et auferentur omnia quae sunt in domo tua et saith the LORD. quae condiderunt patres tui usque in diem hanc in Babylone non remanebit quicquam ait Dominus 20:18 sed et de filiis 20:18 And of thy sons that shall issue from thee, which

thou shalt beget, shall they take away; and they shall be eunuchs in the palace of the king of Babylon.

tuis qui egredientur ex te quos generabis tollentur et erunt eunuchi in palatio regis Babylonis 20:19 dixit Ezechias ad 20:19 Then said Hezekiah unto Isaiah, Good is the word

of the LORD which thou hast spoken. And he said, Is it not good, if peace and truth be in my days?

Esaiam bonus sermo Domini quem locutus est sit pax et ueritas in diebus meis 20:20 reliqua autem sermonum Eze- 20:20 And the rest of the acts of Hezekiah, and all his

might, and how he made a pool, and a conduit, and brought water into the city, are they not written in the book of the

chiae et omnis fortitudo eius et quomodo fecerit piscinam et chronicles of the kings of Judah? aquaeductum et introduxerit aquas in ciuitatem nonne haec

scripta sunt in libro sermonum dierum regum Iuda 20:21 20:21 And Hezekiah slept with his fathers: and Manasseh

his son reigned in his stead.

dormiuitque Ezechias cum patribus suis et regnauit Manasses filius eius pro eo

21:1 duodecim annorum erat Manasses cum regnare coe- 21:1 Manasseh was twelve years old when he began to

reign, and reigned fifty and five years in Jerusalem. And his mother’s name was Hephzibah.

pisset et quinquaginta quinque annis regnauit in Hierusalem nomen matris eius Aphsiba 21:2 fecitque malum in con- 21:2 And he did that which was evil in the sight of the

LORD, after the abominations of the heathen, whom the LORD cast out before the children of Israel.

spectu Domini iuxta idola gentium quas deleuit Dominus a facie filiorum Israhel 21:3 conuersusque est et aedificauit 21:3 For he built up again the high places which Hezekiah

his father had destroyed; and he reared up altars for Baal, and made a grove, as did Ahab king of Israel; and wor-

excelsa quae dissipauerat Ezechias pater eius et erexit aras shipped all the host of heaven, and served them. Baal et fecit lucos sicut fecerat Ahab rex Israhel et adorauit omnem militiam caeli et coluit eam 21:4 extruxitque aras 21:4 And he built altars in the house of the LORD, of

which the LORD said, In Jerusalem will I put my name.

in domo Domini de qua dixit Dominus in Hierusalem ponam nomen meum 21:5 et extruxit altaria uniuersae mili- 21:5 And he built altars for all the host of heaven in the

two courts of the house of the LORD.

tiae caeli in duobus atriis templi Domini 21:6 et transduxit 21:6 And he made his son pass through the fire, and ob-

served times, and used enchantments, and dealt with familiar spirits and wizards: he wrought much wickedness

filium suum per ignem et ariolatus est et obseruauit auguria in the sight of the LORD, to provoke him to anger. et fecit pythones et aruspices multiplicauit ut faceret malum coram Domino et inritaret eum 21:7 posuit quoque idolum 21:7 And he set a graven image of the grove that he had

made in the house, of which the LORD said to David, and to Solomon his son, In this house, and in Jerusalem, which I have chosen out of all tribes of Israel, will I put my name for ever:

luci quem fecerat in templo Domini super quo locutus est Dominus ad Dauid et ad Salomonem filium eius in templo hoc et in Hierusalem quam elegi de cunctis tribubus Israhel

Biblia Sacra Vulgata

21:8 Neither will I make the feet of Israel move any more ponam nomen meum in sempiternum 21:8 et ultra non faout of the land which I gave their fathers; only if they will observe to do according to all that I have commanded them, and according to all the law that my servant Moses commanded them.

ciam commoueri pedem Israhel de terra quam dedi patribus eorum sic tamen si custodierint opere omnia quae praecepi eis et uniuersam legem quam mandauit eis seruus meus Mo-

21:9 But they hearkened not: and Manasseh seduced them ses 21:9 illi uero non audierunt sed seducti sunt a Manasse

to do more evil than did the nations whom the LORD destroyed before the children of Israel.

ut facerent malum super gentes quas contriuit Dominus a fa-

21:10 And the LORD spake by his servants the prophets, cie filiorum Israhel 21:10 locutusque est Dominus in manu

saying,

21:11 Because Manasseh king of Judah hath done these seruorum suorum prophetarum dicens 21:11 quia fecit Ma-

abominations, and hath done wickedly above all that the Amorites did, which were before him, and hath made Ju-

dah also to sin with his idols: nasses rex Iuda abominationes istas pessimas super omnia

quae fecerunt Amorrei ante eum et peccare fecit etiam Iu-

21:12 Therefore thus saith the LORD God of Israel, Be- dam in inmunditiis suis 21:12 propterea haec dicit Domi-

hold, I am bringing such evil upon Jerusalem and Judah, that whosoever heareth of it, both his ears shall tingle.

nus Deus Israhel ecce ego inducam mala super Hierusalem et Iudam ut quicumque audierit tinniant ambae aures eius

21:13 And I will stretch over Jerusalem the line of 21:13 et extendam super Hierusalem funiculum Samariae et

Samaria, and the plummet of the house of Ahab: and I will wipe Jerusalem as a man wipeth a dish, wiping it, and

turning it upside down. pondus domus Ahab et delebo Hierusalem sicut deleri solent

tabulae delens uertam et ducam crebrius stilum super faciem

21:14 And I will forsake the remnant of mine inheritance, eius 21:14 dimittam uero reliquias hereditatis meae et tra-

and deliver them into the hand of their enemies; and they shall become a prey and a spoil to all their enemies;

dam eas in manu inimicorum eius eruntque in uastitate et ra-

21:15 Because they have done that which was evil in my pina cunctis aduersariis suis 21:15 eo quod fecerint malum

sight, and have provoked me to anger, since the day their fathers came forth out of Egypt, even unto this day.

coram me et perseuerauerint inritantes me ex die qua egressi

21:16 Moreover Manasseh shed innocent blood very sunt patres eorum ex Aegypto usque ad diem hanc 21:16 in-

much, till he had filled Jerusalem from one end to another; beside his sin wherewith he made Judah to sin, in doing

that which was evil in the sight of the LORD. super et sanguinem innoxium fudit Manasses multum nimis donec impleret Hierusalem usque ad os absque peccatis suis quibus peccare fecit Iudam ut faceret malum coram Domino

21:17 Now the rest of the acts of Manasseh, and all that he 21:17 reliqua autem sermonum Manasse et uniuersa quae fe-

did, and his sin that he sinned, are they not written in the book of the chronicles of the kings of Judah?

cit et peccatum eius quod peccauit nonne haec scripta sunt

21:18 And Manasseh slept with his fathers, and was buried in libro sermonum dierum regum Iuda 21:18 dormiuitque in the garden of his own house, in the garden of Uzza: and Amon his son reigned in his stead.

Manasses cum patribus suis et sepultus est in horto domus

21:19 Amon was twenty and two years old when he began suae in horto Aza et regnauit Amon filius eius pro eo 21:19

to reign, and he reigned two years in Jerusalem. And his mother’s name was Meshullemeth, the daughter of Haruz

of Jotbah. uiginti et duo annorum erat Amon cum regnare coepisset

duobusque annis regnauit in Hierusalem nomen matris eius

21:20 And he did that which was evil in the sight of the Mesallemeth filia Arus de Iethba 21:20 fecitque malum in

LORD, as his father Manasseh did.

21:21 And he walked in all the way that his father walked conspectu Domini sicut fecerat Manasses pater eius 21:21

in, and served the idols that his father served, and worshipped them:

Liber II Regum

et ambulauit in omni uia per quam ambulauerat pater eius seruiuitque inmunditiis quibus seruierat pater suus et adorauit eas 21:22 et dereliquit Dominum Deum patrum suorum 21:22 And he forsook the LORD God of his fathers, and

walked not in the way of the LORD.

et non ambulauit in uia Domini 21:23 tetenderuntque ei in- 21:23 And the servants of Amon conspired against him,

and slew the king in his own house.

sidias serui sui et interfecerunt regem in domo sua 21:24 21:24 And the people of the land slew all them that had

conspired against king Amon; and the people of the land made Josiah his son king in his stead.

percussit autem populus terrae omnes qui coniurauerant contra regem Amon et constituerunt sibi regem Iosiam filium eius pro eo 21:25 reliqua autem sermonum Amon quae fe- 21:25 Now the rest of the acts of Amon which he did, are

they not written in the book of the chronicles of the kings of Judah?

cit nonne haec scripta sunt in libro sermonum dierum regum Iuda 21:26 sepelieruntque eum in sepulchro suo in horto 21:26 And he was buried in his sepulchre in the garden of

Uzza: and Josiah his son reigned in his stead.

Aza et regnauit Iosias filius eius pro eo

22:1 octo annorum erat Iosias cum regnare coepisset et tri- 22:1 Josiah was eight years old when he began to reign,

and he reigned thirty and one years in Jerusalem. And his mother’s name was Jedidah, the daughter of Adaiah of

ginta uno anno regnauit in Hierusalem nomen matris eius Boscath. Idida filia Phadaia de Besecath 22:2 fecitque quod placitum 22:2 And he did that which was right in the sight of the

LORD, and walked in all the way of David his father, and turned not aside to the right hand or to the left.

erat coram Domino et ambulauit per omnes uias Dauid patris sui non declinauit ad dextram siue ad sinistram 22:3 anno 22:3 And it came to pass in the eighteenth year of king

Josiah, that the king sent Shaphan the son of Azaliah, the son of Meshullam, the scribe, to the house of the LORD,

autem octauodecimo regis Iosiae misit rex Saphan filium saying, Aslia filii Mesullam scribam templi Domini dicens ei 22:4 22:4 Go up to Hilkiah the high priest, that he may sum the

silver which is brought into the house of the LORD, which the keepers of the door have gathered of the people:

uade ad Helciam sacerdotem magnum ut confletur pecunia quae inlata est in templum Domini quam collegerunt ianitores a populo 22:5 deturque fabris per praepositos in domo 22:5 And let them deliver it into the hand of the doers

of the work, that have the oversight of the house of the LORD: and let them give it to the doers of the work which is in the house of the LORD, to repair the breaches of the house,

Domini qui et distribuent eam his qui operantur in templo Domini ad instauranda sarta tecta templi 22:6 tignariis ui- 22:6 Unto carpenters, and builders, and masons, and to

buy timber and hewn stone to repair the house.

delicet et cementariis et his qui interrupta conponunt et ut emantur ligna et lapides de lapidicinis ad instaurandum templum 22:7 uerumtamen non supputetur eis argentum quod 22:7 Howbeit there was no reckoning made with them of

the money that was delivered into their hand, because they dealt faithfully.

accipiunt sed in potestate habeant et in fide 22:8 dixit autem

22:8 And Hilkiah the high priest said unto Shaphan the scribe, I have found the book of the law in the house of the LORD. And Hilkiah gave the book to Shaphan, and he read it.

Helcias pontifex ad Saphan scribam librum legis repperi in domo Domini deditque Helcias uolumen Saphan qui et legit illud 22:9 uenit quoque Saphan scriba ad regem et renuntia- 22:9 And Shaphan the scribe came to the king, and

brought the king word again, and said, Thy servants have gathered the money that was found in the house, and have delivered it into the hand of them that do the work, that have the oversight of the house of the LORD.

uit ei quod praeceperat et ait conflauerunt serui tui pecuniam quae repperta est in domo Domini et dederunt ut distribueretur fabris a praefectis operum templi Domini 22:10 narra- 22:10 And Shaphan the scribe shewed the king, saying,

Hilkiah the priest hath delivered me a book. And Shaphan read it before the king.

Biblia Sacra Vulgata

uitque Saphan scriba regi dicens librum dedit mihi Helcias

22:11 And it came to pass, when the king had heard the sacerdos quem cum legisset Saphan coram rege 22:11 et

words of the book of the law, that he rent his clothes.

audisset rex uerba libri legis Domini scidit uestimenta sua

22:12 And the king commanded Hilkiah the priest, and 22:12 et praecepit Helciae sacerdoti et Ahicham filio Saphan

Ahikam the son of Shaphan, and Achbor the son of Michaiah, and Shaphan the scribe, and Asahiah a servant of the

king’s, saying, et Achobor filio Micha et Saphan scribae et Asaiae seruo 22:13 Go ye, enquire of the LORD for me, and for the peo- regis dicens 22:13 ite et consulite Dominum super me et

ple, and for all Judah, concerning the words of this book that is found: for great is the wrath of the LORD that is kindled against us, because our fathers have not hearkened unto the words of this book, to do according unto all that which is written concerning us.

super populo et super omni Iuda de uerbis uoluminis istius quod inuentum est magna enim ira Domini succensa est contra nos quia non audierunt patres nostri uerba libri huius ut

22:14 So Hilkiah the priest, and Ahikam, and Achbor, and facerent omne quod scriptum est nobis 22:14 ierunt itaque

Shaphan, and Asahiah, went unto Huldah the prophetess, the wife of Shallum the son of Tikvah, the son of Harhas, keeper of the wardrobe; (now she dwelt in Jerusalem in the college;) and they communed with her.

Helcias sacerdos et Ahicham et Achobor et Saphan et Asaia ad Oldam propheten uxorem Sellum filii Thecue filii Araas custodis uestium quae habitabat in Hierusalem in secunda

22:15 And she said unto them, Thus saith the LORD God locutique sunt ad eam 22:15 et illa respondit eis haec di-

of Israel, Tell the man that sent you to me,

cit Dominus Deus Israhel dicite uiro qui misit uos ad me

22:16 Thus saith the LORD, Behold, I will bring evil upon 22:16 haec dicit Dominus ecce ego adducam mala super lothis place, and upon the inhabitants thereof, even all the words of the book which the king of Judah hath read:

cum hunc et super habitatores eius omnia uerba legis quae

22:17 Because they have forsaken me, and have burned in- legit rex Iuda 22:17 quia dereliquerunt me et sacrificaue-

cense unto other gods, that they might provoke me to anger with all the works of their hands; therefore my wrath shall

be kindled against this place, and shall not be quenched. runt diis alienis inritantes me in cunctis operibus manuum

suarum et succendetur indignatio mea in loco hoc et non ex-

22:18 But to the king of Judah which sent you to enquire tinguetur 22:18 regi autem Iuda qui misit uos ut consule-

of the LORD, thus shall ye say to him, Thus saith the LORD God of Israel, As touching the words which thou

hast heard; retis Dominum sic dicetis haec dicit Dominus Deus Israhel 22:19 Because thine heart was tender, and thou hast hum- pro eo quod audisti uerba uoluminis 22:19 et perterritum est

bled thyself before the LORD, when thou heardest what I spake against this place, and against the inhabitants thereof, that they should become a desolation and a curse, and hast rent thy clothes, and wept before me; I also have heard thee, saith the LORD.

cor tuum et humiliatus es coram Domino auditis sermonibus contra locum istum et habitatores eius quo uidelicet fierent in stuporem et in maledictum et scidisti uestimenta tua et

22:20 Behold therefore, I will gather thee unto thy fathers, fleuisti coram me et ego audiui ait Dominus 22:20 idcirco and thou shalt be gathered into thy grave in peace; and thine eyes shall not see all the evil which I will bring upon

this place. And they brought the king word again. colligam te ad patres tuos et colligeris ad sepulchrum tuum in pace ut non uideant oculi tui omnia mala quae inducturus sum super locum istum

23:1 And the king sent, and they gathered unto him all the 23:1 et renuntiauerunt regi quod dixerat qui misit et con-

elders of Judah and of Jerusalem.

23:2 And the king went up into the house of the LORD, gregati sunt ad eum omnes senes Iuda et Hierusalem 23:2

and all the men of Judah and all the inhabitants of Jerusalem with him, and the priests, and the prophets, and all the people, both small and great: and he read in their ears all the words of the book of the covenant which was found in the house of the LORD.

ascenditque rex templum Domini et omnes uiri Iuda uniuer-

Liber II Regum

sique qui habitant in Hierusalem cum eo sacerdotes et prophetae et omnis populus a paruo usque ad magnum legitque cunctis audientibus omnia uerba libri foederis qui inuentus est in domo Domini 23:3 stetitque rex super gradum 23:3 And the king stood by a pillar, and made a covenant

before the LORD, to walk after the LORD, and to keep his commandments and his testimonies and his statutes with all their heart and all their soul, to perform the words of this covenant that were written in this book. And all the people stood to the covenant.

et percussit foedus coram Domino ut ambularent post Dominum et custodirent praecepta eius et testimonia et caerimonias in omni corde et in tota anima et suscitarent uerba foederis huius quae scripta erant in libro illo adquieuitque populus pacto 23:4 et praecepit rex Helciae pontifici et sacer- 23:4 And the king commanded Hilkiah the high priest, and

the priests of the second order, and the keepers of the door, to bring forth out of the temple of the LORD all the vessels that were made for Baal, and for the grove, and for all the host of heaven: and he burned them without Jerusalem in the fields of Kidron, and carried the ashes of them unto Bethel.

dotibus secundi ordinis et ianitoribus ut proicerent de templo Domini omnia uasa quae facta fuerant Baal et in luco et uniuersae militiae caeli et conbusit ea foris Hierusalem in conualle Cedron et tulit puluerem eorum in Bethel 23:5 et de- 23:5 And he put down the idolatrous priests, whom the

kings of Judah had ordained to burn incense in the high places in the cities of Judah, and in the places round about Jerusalem; them also that burned incense unto Baal, to the sun, and to the moon, and to the planets, and to all the host of heaven.

leuit aruspices quos posuerant reges Iuda ad sacrificandum in excelsis per ciuitates Iuda et in circuitu Hierusalem et eos qui adolebant incensum Baal et soli et lunae et duodecim signis et omni militiae caeli 23:6 et efferri fecit lucum de 23:6 And he brought out the grove from the house of the

LORD, without Jerusalem, unto the brook Kidron, and burned it at the brook Kidron, and stamped it small to powder, and cast the powder thereof upon the graves of the children of the people.

domo Domini foras Hierusalem in conualle Cedron et conbusit eum ibi et redeeit in puluerem et proiecit super sepulchrum uulgi 23:7 destruxit quoque aediculas effeminatorum 23:7 And he brake down the houses of the sodomites, that

were by the house of the LORD, where the women wove hangings for the grove.

quae erant in domo Domini pro quibus mulieres texebant quasi domunculas luci 23:8 congregauitque omnes sacer- 23:8 And he brought all the priests out of the cities of

Judah, and defiled the high places where the priests had burned incense, from Geba to Beersheba, and brake down the high places of the gates that were in the entering in of the gate of Joshua the governor of the city, which were on a man’s left hand at the gate of the city.

dotes de ciuitatibus Iuda et contaminauit excelsa ubi sacrificabant sacerdotes de Gabaa usque Bersabee et destruxit aras portarum in introitu ostii Iosue principis ciuitatis quod erat ad sinistram portae ciuitatis 23:9 uerumtamen non ascen- 23:9 Nevertheless the priests of the high places came not

up to the altar of the LORD in Jerusalem, but they did eat of the unleavened bread among their brethren.

debant sacerdotes excelsorum ad altare Domini in Hierusalem sed tantum comedebant azyma in medio fratrum suorum 23:10 contaminauit quoque Thafeth quod est in con- 23:10 And he defiled Topheth, which is in the valley of

the children of Hinnom, that no man might make his son or his daughter to pass through the fire to Molech.

ualle filii Ennom ut nemo consecraret filium suum aut filiam per ignem Moloch 23:11 abstulit quoque equos quos dede- 23:11 And he took away the horses that the kings of Judah

had given to the sun, at the entering in of the house of the LORD, by the chamber of Nathanmelech the chamberlain, which was in the suburbs, and burned the chariots of the sun with fire.

rant reges Iudae soli in introitu templi Domini iuxta exedram Nathanmelech eunuchi qui erat in Farurim currus autem so-

Biblia Sacra Vulgata

23:12 And the altars that were on the top of the upper lis conbusit igni 23:12 altaria quoque quae erant super tecta

chamber of Ahaz, which the kings of Judah had made, and the altars which Manasseh had made in the two courts of the house of the LORD, did the king beat down, and brake them down from thence, and cast the dust of them into the brook Kidron.

cenaculi Ahaz quae fecerant reges Iuda et altaria quae fecerat Manasses in duobus atriis templi Domini destruxit rex et cucurrit inde et dispersit cinerem eorum in torrentem Ced-

23:13 And the high places that were before Jerusalem, ron 23:13 excelsa quoque quae erant in Hierusalem ad de-

which were on the right hand of the mount of corruption, which Solomon the king of Israel had builded for Ashtoreth the abomination of the Zidonians, and for Chemosh the abomination of the Moabites, and for Milcom the abomination of the children of Ammon, did the king defile.

xteram partem montis Offensionis quae aedificauerat Salomon rex Israhel Astharoth idolo Sidoniorum et Chamos offensioni Moab et Melchom abominationi filiorum Ammon

23:14 And he brake in pieces the images, and cut down polluit rex 23:14 et contriuit statuas et succidit lucos reple-

the groves, and filled their places with the bones of men.

23:15 Moreover the altar that was at Bethel, and the high uitque loca eorum ossibus mortuorum 23:15 insuper et al-

place which Jeroboam the son of Nebat, who made Israel to sin, had made, both that altar and the high place he brake down, and burned the high place, and stamped it small to powder, and burned the grove.

tare quod erat in Bethel excelsum quod fecerat Hieroboam filius Nabath qui peccare fecit Israhel et altare illud et excelsum destruxit atque conbusit et comminuit in puluerem suc-

23:16 And as Josiah turned himself, he spied the sepul- cenditque etiam lucum 23:16 et conuersus Iosias uidit ibi

chres that were there in the mount, and sent, and took the bones out of the sepulchres, and burned them upon the altar, and polluted it, according to the word of the LORD which the man of God proclaimed, who proclaimed these words.

sepulchra quae erant in monte misitque et tulit ossa de sepulchris et conbusit ea super altare et polluit illud iuxta uerbum Domini quod locutus est uir Dei qui praedixerat uerba haec

23:17 Then he said, What title is that that I see? And the 23:17 et ait quis est titulus ille quem uideo responderuntque

men of the city told him, It is the sepulchre of the man of God, which came from Judah, and proclaimed these things

that thou hast done against the altar of Bethel. ei ciues illius urbis sepulchrum est hominis Dei qui uenit de

Iuda et praedixit uerba haec quae fecisti super altare Bet-

23:18 And he said, Let him alone; let no man move his hel 23:18 et ait dimittite eum nemo commoueat ossa eius

bones. So they let his bones alone, with the bones of the prophet that came out of Samaria.

et intacta manserunt ossa illius cum ossibus prophetae qui

23:19 And all the houses also of the high places that were uenerat de Samaria 23:19 insuper et omnia fana excelsoin the cities of Samaria, which the kings of Israel had made to provoke the Lord to anger, Josiah took away, and did to

them according to all the acts that he had done in Bethel. rum quae erant in ciuitatibus Samariae quae fecerant reges

Israhel ad inritandum Dominum abstulit Iosias et fecit eis se-

23:20 And he slew all the priests of the high places that cundum omnia opera quae fecerat in Bethel 23:20 et occidit

were there upon the altars, and burned men’s bones upon them, and returned to Jerusalem.

uniuersos sacerdotes excelsorum qui erant ibi super altaria et conbusit ossa humana super ea reuersusque est Hierusa-

23:21 And the king commanded all the people, saying, lem 23:21 et praecepit omni populo dicens facite phase Do-

Keep the passover unto the LORD your God, as it is written in the book of this covenant.

mino Deo uestro secundum quod scriptum est in libro foede-

23:22 Surely there was not holden such a passover from ris huius 23:22 nec enim factum est phase tale a diebus iu-

the days of the judges that judged Israel, nor in all the days of the kings of Israel, nor of the kings of Judah;

dicum qui iudicauerunt Israhel et omnium dierum regum Is-

23:23 But in the eighteenth year of king Josiah, wherein rahel et regum Iuda 23:23 sicut in octauodecimo anno regis

this passover was holden to the LORD in Jerusalem.

Liber II Regum

Iosiae factum est phase istud Domino in Hierusalem 23:24 23:24 Moreover the workers with familiar spirits, and

the wizards, and the images, and the idols, and all the abominations that were spied in the land of Judah and in Jerusalem, did Josiah put away, that he might perform the words of the law which were written in the book that Hilkiah the priest found in the house of the LORD.

sed et pythones et ariolos et figuras idolorum et inmunditias abominationesque quae fuerant in terra Iuda et in Hierusalem abstulit Iosias ut statueret uerba legis quae scripta sunt in libro quem inuenit Helcias sacerdos in templo Domini 23:25 similis illi non fuit ante eum rex qui reuerteretur 23:25 And like unto him was there no king before him,

that turned to the LORD with all his heart, and with all his soul, and with all his might, according to all the law of

ad Dominum in omni corde suo et in tota anima sua et in uni- Moses; neither after him arose there any like him. uersa uirtute sua iuxta omnem legem Mosi neque post eum surrexit similis illi 23:26 uerumtamen non est auersus Do- 23:26 Notwithstanding the LORD turned not from the

fierceness of his great wrath, wherewith his anger was kindled against Judah, because of all the provocations that

minus ab ira furoris sui magni quo iratus est furor eius contra Manasseh had provoked him withal. Iudam propter inritationes quibus prouocauerat eum Manasses 23:27 dixit itaque Dominus etiam Iudam auferam a fa- 23:27 And the LORD said, I will remove Judah also out

of my sight, as I have removed Israel, and will cast off this city Jerusalem which I have chosen, and the house of

cie mea sicut abstuli Israhel et proiciam ciuitatem hanc quam which I said, My name shall be there. elegi Hierusalem et domum de qua dixi erit nomen meum ibi 23:28 reliqua autem uerba Iosiae et uniuersa quae fe- 23:28 Now the rest of the acts of Josiah, and all that he

did, are they not written in the book of the chronicles of the kings of Judah?

cit nonne haec scripta sunt in libro uerborum dierum regum Iuda 23:29 in diebus eius ascendit Pharao Necho rex Ae- 23:29 In his days Pharaohnechoh king of Egypt went up

against the king of Assyria to the river Euphrates: and king Josiah went against him; and he slew him at Megiddo,

gypti contra regem Assyriorum ad flumen Eufraten et abiit when he had seen him. Iosias rex in occursum eius et occisus est in Mageddo cum uidisset eum 23:30 et portauerunt eum serui sui mortuum 23:30 And his servants carried him in a chariot dead from

Megiddo, and brought him to Jerusalem, and buried him in his own sepulchre. And the people of the land took Jehoahaz the son of Josiah, and anointed him, and made him king in his father’s stead.

de Mageddo et pertulerunt in Hierusalem et sepelierunt eum in sepulchro suo tulitque populus terrae Ioahaz filium Iosiae et unxerunt eum et constituerunt eum regem pro patre suo 23:31 uiginti trium annorum erat Ioahaz cum regnare 23:31 Jehoahaz was twenty and three years old when he

began to reign; and he reigned three months in Jerusalem. And his mother’s name was Hamutal, the daughter of

coepisset et tribus mensibus regnauit in Hierusalem nomen Jeremiah of Libnah. matris eius Amithal filia Hieremiae de Lobna 23:32 et fe- 23:32 And he did that which was evil in the sight of the

LORD, according to all that his fathers had done.

cit malum coram Domino iuxta omnia quae fecerant patres eius 23:33 uinxitque eum Pharao Necho in Rebla quae est 23:33 And Pharaohnechoh put him in bands at Riblah in

the land of Hamath, that he might not reign in Jerusalem; and put the land to a tribute of an hundred talents of silver,

in terra Emath ne regnaret in Hierusalem et inposuit mul- and a talent of gold. tam terrae centum talentis argenti et talento auri 23:34 re- 23:34 And Pharaohnechoh made Eliakim the son of Josiah

king in the room of Josiah his father, and turned his name to Jehoiakim, and took Jehoahaz away: and he came to

gemque constituit Pharao Necho Eliachim filium Iosiae pro Egypt, and died there. Iosia patre eius uertitque nomen eius Ioiachim porro Ioahaz tulit et duxit in Aegyptum 23:35 argentum autem et aurum 23:35 And Jehoiakim gave the silver and the gold to

Pharaoh; but he taxed the land to give the money according to the commandment of Pharaoh: he exacted the silver and the gold of the people of the land, of every one according to his taxation, to give it unto Pharaohnechoh.

Biblia Sacra Vulgata

dedit Ioiachim Pharaoni cum indixisset terrae per singulos ut conferretur iuxta praeceptum Pharaonis et unumquemque secundum uires suas exegit tam argentum quam aurum de

23:36 Jehoiakim was twenty and five years old when he populo terrae ut daret Pharaoni Necho 23:36 uiginti quin-

began to reign; and he reigned eleven years in Jerusalem. And his mother’s name was Zebudah, the daughter of

Pedaiah of Rumah. que annorum erat Ioiachim cum regnare coepisset et unde-

cim annis regnauit in Hierusalem nomen matris eius Zebida

23:37 And he did that which was evil in the sight of the filia Phadaia de Ruma 23:37 et fecit malum coram Domino

LORD, according to all that his fathers had done.

iuxta omnia quae fecerant patres eius

24:1 In his days Nebuchadnezzar king of Babylon came 24:1 in diebus eius ascendit Nabuchodonosor rex Baby-

up, and Jehoiakim became his servant three years: then he turned and rebelled against him.

lonis et factus est ei Ioiachim seruus tribus annis et rursum

24:2 And the LORD sent against him bands of the rebellauit contra eum 24:2 inmisitque ei Dominus latruncu-

Chaldees, and bands of the Syrians, and bands of the Moabites, and bands of the children of Ammon, and sent them against Judah to destroy it, according to the word of the LORD, which he spake by his servants the prophets.

los Chaldeorum et latrunculos Syriae latrunculos Moab et latrunculos filiorum Ammon et inmisit eos in Iudam ut disperderent eum iuxta uerbum Domini quod locutus erat per ser-

24:3 Surely at the commandment of the LORD came this uos suos prophetas 24:3 factum est autem hoc per uerbum

upon Judah, to remove them out of his sight, for the sins of Manasseh, according to all that he did;

Domini contra Iudam ut auferret eum coram se propter pec-

24:4 And also for the innocent blood that he shed: for cata Manasse uniuersa quae fecit 24:4 et propter sanguinem

he filled Jerusalem with innocent blood; which the LORD would not pardon.

innoxium quem effudit et impleuit Hierusalem cruore inno-

24:5 Now the rest of the acts of Jehoiakim, and all that he centium et ob hanc rem noluit Dominus propitiari 24:5 reli-

did, are they not written in the book of the chronicles of the kings of Judah?

qua autem sermonum Ioiachim et uniuersa quae fecit nonne haec scripta sunt in libro sermonum dierum regum Iuda et

24:6 So Jehoiakim slept with his fathers: and Jehoiachin dormiuit Ioiachim cum patribus suis 24:6 regnauitque Io-

his son reigned in his stead.

24:7 And the king of Egypt came not again any more out iachin filius eius pro eo 24:7 et ultra non addidit rex Ae-

of his land: for the king of Babylon had taken from the river of Egypt unto the river Euphrates all that pertained to

the king of Egypt. gypti ut egrederetur de terra sua tulerat enim rex Babylonis

a riuo Aegypti usque ad fluuium Eufraten omnia quae fue-

24:8 Jehoiachin was eighteen years old when he began to rant regis Aegypti 24:8 decem et octo annorum erat Ioia-

reign, and he reigned in Jerusalem three months. And his mother’s name was Nehushta, the daughter of Elnathan of

Jerusalem. chin cum regnare coepisset et tribus mensibus regnauit in

Hierusalem nomen matris eius Naestha filia Helnathan de

24:9 And he did that which was evil in the sight of the Hierusalem 24:9 et fecit malum coram Domino iuxta om-

LORD, according to all that his father had done.

24:10 At that time the servants of Nebuchadnezzar king nia quae fecerat pater eius 24:10 in tempore illo ascende-

of Babylon came up against Jerusalem, and the city was besieged.

runt serui Nabuchodonosor regis Babylonis in Hierusalem

24:11 And Nebuchadnezzar king of Babylon came against et circumdata est urbs munitionibus 24:11 uenitque Nabu-

the city, and his servants did besiege it.

chodonosor rex Babylonis ad ciuitatem cum serui eius ob-

Liber II Regum

pugnarent eam 24:12 egressusque est Ioiachin rex Iuda ad 24:12 And Jehoiachin the king of Judah went out to the

king of Babylon, he, and his mother, and his servants, and his princes, and his officers: and the king of Babylon took

regem Babylonis ipse et mater eius et serui eius et princi- him in the eighth year of his reign. pes eius et eunuchi eius et suscepit eum rex Babylonis anno octauo regni sui 24:13 et protulit inde omnes thesauros do- 24:13 And he carried out thence all the treasures of the

house of the LORD, and the treasures of the king’s house, and cut in pieces all the vessels of gold which Solomon king of Israel had made in the temple of the LORD, as the LORD had said.

mus Domini et thesauros domus regiae et concidit uniuersa uasa aurea quae fecerat Salomon rex Israhel in templo Domini iuxta uerbum Domini 24:14 et transtulit omnem Hie- 24:14 And he carried away all Jerusalem, and all the

princes, and all the mighty men of valour, even ten thousand captives, and all the craftsmen and smiths: none re-

rusalem et uniuersos principes et omnes fortes exercitus de- mained, save the poorest sort of the people of the land. cem milia in captiuitatem et omnem artificem et clusorem nihilque relictum est exceptis pauperibus populi terrae 24:15 24:15 And he carried away Jehoiachin to Babylon, and the

king’s mother, and the king’s wives, and his officers, and the mighty of the land, those carried he into captivity from

transtulit quoque Ioiachin in Babylonem et matrem regis et Jerusalem to Babylon. uxores regis et eunuchos eius et iudices terrae duxit in captiuitatem de Hierusalem in Babylonem 24:16 et omnes uiros 24:16 And all the men of might, even seven thousand, and

craftsmen and smiths a thousand, all that were strong and apt for war, even them the king of Babylon brought captive

robustos septem milia et artifices et clusores mille omnes ui- to Babylon. ros fortes et bellatores duxitque eos rex Babylonis captiuos in Babylonem 24:17 et constituit Matthaniam patruum eius 24:17 And the king of Babylon made Mattaniah his fa-

ther’s brother king in his stead, and changed his name to Zedekiah.

pro eo inposuitque nomen ei Sedeciam 24:18 uicesimum et

24:18 Zedekiah was twenty and one years old when he began to reign, and he reigned eleven years in Jerusalem. And his mother’s name was Hamutal, the daughter of Jeremiah of Libnah.

primum annum aetatis habebat Sedecias cum regnare coepisset et undecim annis regnauit in Hierusalem nomen matris eius erat Amithal filia Hieremiae de Lobna 24:19 et fe- 24:19 And he did that which was evil in the sight of the

LORD, according to all that Jehoiakim had done.

cit malum coram Domino iuxta omnia quae fecerat Ioiachim 24:20 irascebatur enim Dominus contra Hierusalem et contra 24:20 For through the anger of the LORD it came to pass

in Jerusalem and Judah, until he had cast them out from his presence, that Zedekiah rebelled against the king of

Iudam donec proiceret eos a facie sua recessitque Sedecias a Babylon. rege Babylonis

25:1 factum est autem anno nono regni eius mense de- 25:1 And it came to pass in the ninth year of his reign, in the tenth month, in the tenth day of the month, that Nebuchadnezzar king of Babylon came, he, and all his host, against Jerusalem, and pitched against it; and they built forts against it round about.

cimo decima die mensis uenit Nabuchodonosor rex Babylonis ipse et omnis exercitus eius in Hierusalem et circumdederunt eam et extruxerunt in circuitu eius munitiones 25:2 25:2 And the city was besieged unto the eleventh year of

king Zedekiah.

et clausa est ciuitas atque uallata usque ad undecimum annum regis Sedeciae 25:3 nona die mensis praeualuitque fa- 25:3 And on the ninth day of the fourth month the famine

prevailed in the city, and there was no bread for the people of the land.

mes in ciuitate nec erat panis populo terrae 25:4 et inter-

25:4 And the city was broken up, and all the men of war fled by night by the way of the gate between two walls, which is by the king’s garden: (now the Chaldees were against the city round about:) and the king went the way toward the plain.

rupta est ciuitas et omnes uiri bellatores nocte fugerunt per uiam portae quae est inter duplicem murum ad hortum re-

Biblia Sacra Vulgata

gis porro Chaldei obsidebant in circuitu ciuitatem fugit ita-

25:5 And the army of the Chaldees pursued after the king, que per uiam quae ducit ad campestria solitudinis 25:5 et

and overtook him in the plains of Jericho: and all his army were scattered from him.

persecutus est exercitus Chaldeorum regem conprehenditque eum in planitie Hiericho et omnes bellatores qui erant

25:6 So they took the king, and brought him up to the king cum eo dispersi sunt et reliquerunt eum 25:6 adprehensum

of Babylon to Riblah; and they gave judgment upon him.

ergo regem duxerunt ad regem Babylonis in Reblatha qui

25:7 And they slew the sons of Zedekiah before his eyes, locutus est cum eo iudicium 25:7 filios autem Sedeciae oc-

and put out the eyes of Zedekiah, and bound him with fetters of brass, and carried him to Babylon.

cidit coram eo et oculos eius effodit uinxitque eum catenis et

25:8 And in the fifth month, on the seventh day of the adduxit in Babylonem 25:8 mense quinto septima die men-

month, which is the nineteenth year of king Nebuchadnezzar king of Babylon, came Nebuzaradan, captain of the

guard, a servant of the king of Babylon, unto Jerusalem: sis ipse est annus nonusdecimus regis Babylonis uenit Na-

buzardan princeps exercitus seruus regis Babylonis Hieru-

25:9 And he burnt the house of the LORD, and the king’s salem 25:9 et succendit domum Domini et domum regis et

house, and all the houses of Jerusalem, and every great man’s house burnt he with fire.

domos Hierusalem omnemque domum conbusit igni 25:10

25:10 And all the army of the Chaldees, that were with the captain of the guard, brake down the walls of Jerusalem round about.

et muros Hierusalem in circuitu destruxit omnis exercitus

25:11 Now the rest of the people that were left in the city, Chaldeorum qui erat cum principe militum 25:11 reliquam and the fugitives that fell away to the king of Babylon, with the remnant of the multitude, did Nebuzaradan the

captain of the guard carry away. autem populi partem qui remanserat in ciuitate et perfugas

qui transfugerant ad regem Babylonis et reliquum uulgus

25:12 But the captain of the guard left of the door of the transtulit Nabuzardan princeps militiae 25:12 et de paupe-

poor of the land to be vinedressers and husbandmen.

25:13 And the pillars of brass that were in the house of the ribus terrae reliquit uinitores et agricolas 25:13 columnas

LORD, and the bases, and the brasen sea that was in the house of the LORD, did the Chaldees break in pieces, and

carried the brass of them to Babylon. autem aereas quae erant in templo Domini et bases et mare

aereum quod erat in domo Domini confregerunt Chaldei et

25:14 And the pots, and the shovels, and the snuffers, and transtulerunt aes omnium in Babylonem 25:14 ollas quothe spoons, and all the vessels of brass wherewith they ministered, took they away.

que aereas et trullas et tridentes et scyphos et omnia uasa

25:15 And the firepans, and the bowls, and such things as aerea in quibus ministrabant tulerunt 25:15 necnon turibula

were of gold, in gold, and of silver, in silver, the captain of the guard took away.

et fialas quae aurea aurea et quae argentea argentea tulit prin-

25:16 The two pillars, one sea, and the bases which ceps militiae 25:16 id est columnas duas mare unum et ba-

Solomon had made for the house of the LORD; the brass of all these vessels was without weight.

ses quas fecerat Salomon in templo Domini non erat pondus

25:17 The height of the one pillar was eighteen cubits, aeris omnium uasorum 25:17 decem et octo cubitos altitu-

and the chapiter upon it was brass: and the height of the chapiter three cubits; and the wreathen work, and pomegranates upon the chapiter round about, all of brass: and like unto these had the second pillar with wreathen work.

dinis habebat columna una et capitellum aereum super se altitudinis trium cubitorum et reticulum et malogranata super capitellum columnae omnia aerea similem et columna

25:18 And the captain of the guard took Seraiah the chief secunda habebat ornatum 25:18 tulit quoque princeps mili-

priest, and Zephaniah the second priest, and the three keepers of the door:

tiae Seraian sacerdotem primum et Sophoniam sacerdotem

Liber II Regum

secundum et tres ianitores 25:19 et de ciuitate eunuchum 25:19 And out of the city he took an officer that was set

over the men of war, and five men of them that were in the king’s presence, which were found in the city, and the principal scribe of the host, which mustered the people of the land, and threescore men of the people of the land that were found in the city:

unum qui erat praefectus super uiros bellatores et quinque uiros de his qui steterant coram rege quos repperit in ciuitate et Sopher principem exercitus qui probabat tirones de populo terrae et sex uiros e uulgo qui inuenti fuerant in ciuitate 25:20 quos tollens Nabuzardan princeps militum duxit 25:20 And Nebuzaradan captain of the guard took these,

and brought them to the king of Babylon to Riblah:

ad regem Babylonis in Reblatha 25:21 percussitque eos rex 25:21 And the king of Babylon smote them, and slew them

at Riblah in the land of Hamath. So Judah was carried away out of their land.

Babylonis et interfecit in Reblatha in terra Emath et translatus est Iuda de terra sua 25:22 populo autem qui relictus erat 25:22 And as for the people that remained in the land of

Judah, whom Nebuchadnezzar king of Babylon had left, even over them he made Gedaliah the son of Ahikam, the

in terra Iuda quem dimiserat Nabuchodonosor rex Babylonis son of Shaphan, ruler. praefecit Godoliam filium Ahicham filii Saphan 25:23 quod 25:23 And when all the captains of the armies, they and

their men, heard that the king of Babylon had made Gedaliah governor, there came to Gedaliah to Mizpah, even Ishmael the son of Nethaniah, and Johanan the son of Careah, and Seraiah the son of Tanhumeth the Netophathite, and Jaazaniah the son of a Maachathite, they and their men.

cum audissent omnes duces militum ipsi et uiri qui erant cum eis uidelicet quod constituisset rex Babylonis Godoliam uenerunt ad Godoliam in Maspha Ismahel filius Nathaniae et Iohanan filius Caree et Sareia filius Thenaameth Nethophathites et Iezonias filius Maachathi ipsi et socii eorum 25:24 25:24 And Gedaliah sware to them, and to their men, and

said unto them, Fear not to be the servants of the Chaldees: dwell in the land, and serve the king of Babylon; and it

iurauitque eis Godolias et sociis eorum dicens nolite timere shall be well with you. seruire Chaldeis manete in terra et seruite regi Babylonis et bene erit uobis 25:25 factum est autem in mense septimo 25:25 But it came to pass in the seventh month, that Ish-

mael the son of Nethaniah, the son of Elishama, of the seed royal, came, and ten men with him, and smote Gedaliah, that he died, and the Jews and the Chaldees that were with him at Mizpah.

uenit Ismahel filius Nathaniae filii Elisama de semine regio et decem uiri cum eo percusseruntque Godoliam qui mortuus est sed et Iudaeos et Chaldeos qui erant cum eo in Maspha 25:26 consurgens autem omnis populus a paruo usque 25:26 And all the people, both small and great, and the

captains of the armies, arose, and came to Egypt: for they were afraid of the Chaldees.

ad magnum et principes militum uenerunt in Aegyptum timentes Chaldeos 25:27 factum est uero anno tricesimo se- 25:27 And it came to pass in the seven and thirtieth year

of the captivity of Jehoiachin king of Judah, in the twelfth month, on the seven and twentieth day of the month, that Evilmerodach king of Babylon in the year that he began to reign did lift up the head of Jehoiachin king of Judah out of prison;

ptimo transmigrationis Ioiachin regis Iudae mense duodecimo uicesima septima die mensis subleuauit Euilmerodach rex Babylonis anno quo regnare coeperat caput Ioiachin regis Iuda de carcere 25:28 et locutus est ei benigna et posuit 25:28 And he spake kindly to him, and set his throne above

the throne of the kings that were with him in Babylon;

thronum eius super thronum regum qui erant cum eo in Babylone 25:29 et mutauit uestes eius quas habuerat in carcere 25:29 And changed his prison garments: and he did eat

bread continually before him all the days of his life.

et comedebat panem semper in conspectu eius cunctis diebus uitae suae 25:30 annonam quoque constituit ei absque 25:30 And his allowance was a continual allowance given

him of the king, a daily rate for every day, all the days of his life.

Biblia Sacra Vulgata

intermissione quae et dabatur ei a rege per singulos dies omnibus diebus uitae suae

Liber I Paralipomenon

1:1 Adam Seth Enos 1:2 Cainan Malelehel Iared 1:3 1:1 Adam, Sheth, Enosh,

1:2 Kenan, Mahalaleel, Jered,

Enoch Matusale Lamech 1:4 Noe Sem Ham et Iafeth 1:5 _{1:3 Henoch, Methuselah, Lamech,}

1:4 Noah, Shem, Ham, and Japheth.

1:5 The sons of Japheth; Gomer, and Magog, and Madai, and Javan, and Tubal, and Meshech, and Tiras.

filii Iafeth Gomer Magog Madai et Iauan Thubal Mosoch Thiras 1:6 porro filii Gomer Aschenez et Rifath et Tho- 1:6 And the sons of Gomer; Ashchenaz, and Riphath, and

Togarmah.

gorma 1:7 filii autem Iauan Elisa et Tharsis Cetthim et Do- 1:7 And the sons of Javan; Elishah, and Tarshish, Kittim,

and Dodanim.

danim 1:8 filii Ham Chus et Mesraim Phut et Chanaan 1:9 1:8 The sons of Ham; Cush, and Mizraim, Put, and

Canaan.

1:9 And the sons of Cush; Seba, and Havilah, and Sabta, and Raamah, and Sabtecha. And the sons of Raamah; Sheba, and Dedan.

filii autem Chus Saba et Euila Sabatha et Rechma et Sabathaca porro filii Rechma Saba et Dadan 1:10 Chus autem ge- 1:10 And Cush begat Nimrod: he began to be mighty upon

the earth.

nuit Nemrod iste coepit esse potens in terra 1:11 Mesraim 1:11 And Mizraim begat Ludim, and Anamim, and

Lehabim, and Naphtuhim,

uero genuit Ludim et Anamim et Laabim et Nepthuim 1:12 1:12 And Pathrusim, and Casluhim, (of whom came the

Philistines,) and Caphthorim.

Phethrosim quoque et Chasluim de quibus egressi sunt Philisthim et Capthurim 1:13 Chanaan uero genuit Sidonem 1:13 And Canaan begat Zidon his firstborn, and Heth, primogenitum et Heth 1:14 Iebuseum quoque et Amorreum 1:14 The Jebusite also, and the Amorite, and the Gir-

gashite,

et Gergeseum 1:15 Euheumque et Aruceum et Asineum 1:15 And the Hivite, and the Arkite, and the Sinite, 1:16 Aradium quoque et Samareum et Ematheum 1:17 fi- 1:16 And the Arvadite, and the Zemarite, and the Hamath-

ite.

1:17 The sons of Shem; Elam, and Asshur, and Arphaxad, and Lud, and Aram, and Uz, and Hul, and Gether, and Meshech.

lii Sem Aelam et Assur et Arfaxad et Lud et Aram et Vs et

Hul et Gothor et Mosoch 1:18 Arfaxad autem genuit Sala 1:18 And Arphaxad begat Shelah, and Shelah begat Eber.

qui et ipse genuit Heber 1:19 porro Heber nati sunt duo filii 1:19 And unto Eber were born two sons: the name of the

one was Peleg; because in his days the earth was divided: and his brother’s name was Joktan.

nomen uni Phaleg quia in diebus eius diuisa est terra et nomen fratris eius Iectan 1:20 Iectan autem genuit Helmodad 1:20 And Joktan begat Almodad, and Sheleph, and Haz-

armaveth, and Jerah,

et Saleph et Asermoth et Iare 1:21 Aduram quoque et Vzal 1:21 Hadoram also, and Uzal, and Diklah, et Decla 1:22 Ebal etiam et Abimahel et Saba necnon 1:23 1:22 And Ebal, and Abimael, and Sheba,

1:23 And Ophir, and Havilah, and Jobab. All these were the sons of Joktan.

et Ophir et Euila et Iobab omnes isti filii Iectan 1:24 Sem

1:24 Shem, Arphaxad, Shelah,

Arfaxad Sale 1:25 Heber Phaleg Raau 1:26 Serug Nahor 1:25 Eber, Peleg, Reu,

1:26 Serug, Nahor, Terah,

Thare 1:27 Abram iste est Abraham 1:28 filii autem Ab- _{1:27 Abram; the same is Abraham.}

1:28 The sons of Abraham; Isaac, and Ishmael.

raham Isaac et Ismahel 1:29 et hae generationes eorum pri- 1:29 These are their generations: The firstborn of Ishmael,

Nebaioth; then Kedar, and Adbeel, and Mibsam,

mogenitus Ismahelis Nabaioth et Cedar et Adbeel et Mabsam 1:30 Masma et Duma Massa Adad et Thema 1:31 Ia- 1:30 Mishma, and Dumah, Massa, Hadad, and Tema,

1:31 Jetur, Naphish, and Kedemah. These are the sons of Ishmael.

thur Naphis Cedma hii sunt filii Ismahelis 1:32 filii autem

1:32 Now the sons of Keturah, Abraham’s concubine: she bare Zimran, and Jokshan, and Medan, and Midian, and Ishbak, and Shuah. And the sons of Jokshan; Sheba, and Dedan.

Cetthurae concubinae Abraham quos genuit Zamram Iecsan Madan Madian Iesboc Sue porro filii Iecsan Saba et Dadan 1:33 filii autem Madian Epha et Apher et Enoch et Abida et 1:33 And the sons of Midian; Ephah, and Epher, and

Henoch, and Abida, and Eldaah. All these are the sons of Keturah.

Eldaa omnes hii filii Cetthurae 1:34 generauit autem Abra-

1:34 And Abraham begat Isaac. The sons of Isaac; Esau and Israel.

1:35 The sons of Esau; Eliphaz, Reuel, and Jeush, and ham Isaac cuius fuerunt filii Esau et Israhel 1:35 filii Esau

Jaalam, and Korah.

1:36 The sons of Eliphaz; Teman, and Omar, Zephi, and Eliphaz Rauhel Iaus Ialam Core 1:36 filii Eliphaz Theman

Gatam, Kenaz, and Timna, and Amalek. 1:37 The sons of Reuel; Nahath, Zerah, Shammah, and Omer Sepphu Gethem Cenez Thamna Amalech 1:37 filii Mizzah. 1:38 And the sons of Seir; Lotan, and Shobal, and Zibeon, Rauhel Naath Zara Samma Maza 1:38 filii Seir Lothan Soand Anah, and Dishon, and Ezar, and Dishan.

1:39 And the sons of Lotan; Hori, and Homam: and Timna bal Sebeon Ana Dison Eser Disan 1:39 filii Lothan Horri

was Lotan’s sister. 1:40 The sons of Shobal; Alian, and Manahath, and Ebal, Humam soror autem Lothan fuit Thamna 1:40 filii Sobal Shephi, and Onam. and the sons of Zibeon; Aiah, and Anah.

Alian et Manaath et Ebal et Sepphi et Onam filii Sebeon

1:41 The sons of Anah; Dishon. And the sons of Dishon; Aia et Ana filii Ana Dison 1:41 filii Dison Amaran et Ese-

Amram, and Eshban, and Ithran, and Cheran.

1:42 The sons of Ezer; Bilhan, and Zavan, and Jakan. The ban et Iethran et Charan 1:42 filii Eser Balaan et Zaban et

sons of Dishan; Uz, and Aran.

1:43 Now these are the kings that reigned in the land of Iacan filii Dison Vs et Aran 1:43 isti sunt reges qui impera-

Edom before any king reigned over the children of Israel; Bela the son of Beor: and the name of his city was Din-

habah. uerunt in terra Edom antequam esset rex super filios Israhel 1:44 And when Bela was dead, Jobab the son of Zerah of Bale filius Beor et nomen ciuitatis eius Denaba 1:44 mor-

Bozrah reigned in his stead.

tuus est autem Bale et regnauit pro eo Iobab filius Zare de

1:45 And when Jobab was dead, Husham of the land of Bosra 1:45 cumque et Iobab fuisset mortuus regnauit pro

the Temanites reigned in his stead.

1:46 And when Husham was dead, Hadad the son of Be- eo Husam de terra Themanorum 1:46 obiit quoque et Hu-

dad, which smote Midian in the field of Moab, reigned in his stead: and the name of his city was Avith.

sam et regnauit pro eo Adad filius Badad qui percussit Ma-

1:47 And when Hadad was dead, Samlah of Masrekah dian in terra Moab et nomen ciuitatis eius Auith 1:47 cum-

reigned in his stead.

que et Adad fuisset mortuus regnauit pro eo Semla de Mas-

1:48 And when Samlah was dead, Shaul of Rehoboth by reca 1:48 sed et Semla mortuus est et regnauit pro eo Saul

the river reigned in his stead.

1:49 And when Shaul was dead, Baalhanan the son of de Rooboth quae iuxta amnem sita est 1:49 mortuo quoque

Achbor reigned in his stead.

1:50 And when Baalhanan was dead, Hadad reigned in his Saul regnauit pro eo Baalanan filius Achobor 1:50 sed et

stead: and the name of his city was Pai; and his wife’s name was Mehetabel, the daughter of Matred, the daugh-

ter of Mezahab. hic mortuus est et regnauit pro eo Adad cuius urbis fuit nomen Phou et appellata est uxor eius Mehetabel filia Matred

1:51 Hadad died also. And the dukes of Edom were; duke filiae Mezaab 1:51 Adad autem mortuo duces pro regibus

Timnah, duke Aliah, duke Jetheth,

in Edom esse coeperunt dux Thamna dux Alua dux Ietheth

1:52 Duke Aholibamah, duke Elah, duke Pinon, 1:52 dux Oolibama dux Hela dux Phinon 1:53 dux Cenez

1:53 Duke Kenaz, duke Teman, duke Mibzar, 1:54 Duke Magdiel, duke Iram. These are the dukes of dux Theman dux Mabsar 1:54 dux Magdihel dux Iram hii Edom.

duces Edom

2:1 These are the sons of Israel; Reuben, Simeon, Levi, 2:1 filii autem Israhel Ruben Symeon Leui Iuda Isachar and Judah, Issachar, and Zebulun, 2:2 Dan, Joseph, and Benjamin, Naphtali, Gad, and Asher. et Zabulon 2:2 Dan Ioseph Beniamin Nepthali Gad Aser

2:3 The sons of Judah; Er, and Onan, and Shelah: which 2:3 filii Iuda Her Aunan Sela tres nati sunt ei de filia Sue

three were born unto him of the daughter of Shua the Canaanitess. And Er, the firstborn of Judah, was evil in

the sight of the LORD; and he slew him. Chananitidis fuit autem Her primogenitus Iuda malus coram 2:4 And Tamar his daughter in law bore him Pharez and Domino et occidit eum 2:4 Thamar autem nurus eius pe-

Zerah. All the sons of Judah were five.

Liber I Paralipomenon

perit ei Phares et Zara omnes ergo filii Iuda quinque 2:5 2:5 The sons of Pharez; Hezron, and Hamul. filii autem Phares Esrom et Hamul 2:6 filii quoque Zarae 2:6 And the sons of Zerah; Zimri, and Ethan, and Heman,

and Calcol, and Dara: five of them in all.

Zamri et Ethan et Eman Chalchal quoque et Darda simul quinque 2:7 filii Carmi Achar qui turbauit Israhel et pec- 2:7 And the sons of Carmi; Achar, the troubler of Israel,

who transgressed in the thing accursed.

cauit in furto anathematis 2:8 filii Ethan Azarias 2:9 filii 2:8 And the sons of Ethan; Azariah.

2:9 The sons also of Hezron, that were born unto him; Jerahmeel, and Ram, and Chelubai.

autem Esrom qui nati sunt ei Ieremahel et Ram et Chalubi 2:10 porro Ram genuit Aminadab Aminadab autem genuit 2:10 And Ram begat Amminadab; and Amminadab begat

Nahshon, prince of the children of Judah;

Naasson principem filiorum Iuda 2:11 Naasson quoque ge- 2:11 And Nahshon begat Salma, and Salma begat Boaz, nuit Salma de quo ortus est Boez 2:12 Boez uero genuit Ob- 2:12 And Boaz begat Obed, and Obed begat Jesse, ed qui et ipse genuit Isai 2:13 Isai autem genuit primogeni- 2:13 And Jesse begat his firstborn Eliab, and Abinadab the

second, and Shimma the third,

tum Heliab secundum Abinadab tertium Samaa 2:14 quar- 2:14 Nethaneel the fourth, Raddai the fifth, tum Nathanahel quintum Raddai 2:15 sextum Asom septi- 2:15 Ozem the sixth, David the seventh: mum Dauid 2:16 quorum sorores fuerunt Saruia et Abigail 2:16 Whose sisters were Zeruiah, and Abigail. And the

sons of Zeruiah; Abishai, and Joab, and Asahel, three.

filii Saruiae Abisai Ioab et Asahel tres 2:17 Abigail autem 2:17 And Abigail bare Amasa: and the father of Amasa

was Jether the Ishmeelite.

genuit Amasa cuius pater fuit Iether Ismahelites 2:18 Cha- 2:18 And Caleb the son of Hezron begat children of

Azubah his wife, and of Jerioth: her sons are these; Jesher, and Shobab, and Ardon.

leb uero filius Esrom accepit uxorem nomine Azuba de qua genuit Ierioth fueruntque filii eius Iesar et Sobab et Ardon 2:19 cumque mortua fuisset Azuba accepit uxorem Chaleb 2:19 And when Azubah was dead, Caleb took unto him

Ephrath, which bare him Hur.

Ephrath quae peperit ei Vr 2:20 porro Vr genuit Vri et Vri 2:20 And Hur begat Uri, and Uri begat Bezaleel. genuit Beselehel 2:21 post haec ingressus est Esrom ad fi- 2:21 And afterward Hezron went in to the daughter of

Machir the father of Gilead, whom he married when he was threescore years old; and she bare him Segub.

liam Machir patris Galaad et accepit eam cum esset annorum sexaginta quae peperit ei Segub 2:22 sed et Segub ge- 2:22 And Segub begat Jair, who had three and twenty

cities in the land of Gilead.

nuit Iair et possedit uiginti tres ciuitates in terra Galaad 2:23 2:23 And he took Geshur, and Aram, with the towns of

Jair, from them, with Kenath, and the towns thereof, even threescore cities. All these belonged to the sons of Machir

cepitque Gessur et Aram oppida Iair et Canath et uiculos the father of Gilead. eius sexaginta ciuitatum omnes isti filii Machir patris Galaad 2:24 cum autem mortuus esset Esrom ingressus est Chaleb ad 2:24 And after that Hezron was dead in Calebephratah,

then Abiah Hezron’s wife bare him Ashur the father of Tekoa.

Ephrata habuit quoque Esrom uxorem Abia quae peperit ei Assur patrem Thecue 2:25 nati sunt autem filii Hieramehel 2:25 And the sons of Jerahmeel the firstborn of Hezron

were, Ram the firstborn, and Bunah, and Oren, and Ozem, and Ahijah.

primogeniti Esrom Ram primogenitus eius et Buna et Aran et Asom et Ahia 2:26 duxit quoque uxorem alteram Hie- 2:26 Jerahmeel had also another wife, whose name was

Atarah; she was the mother of Onam.

ramehel nomine Atara quae fuit mater Onam 2:27 sed et 2:27 And the sons of Ram the firstborn of Jerahmeel were,

Maaz, and Jamin, and Eker.

filii Ram primogeniti Hieramehel fuerunt Moos et Iamin et Achar 2:28 Onam autem habuit filios Semmei et Iada fi- 2:28 And the sons of Onam were, Shammai, and Jada.

And the sons of Shammai; Nadab and Abishur.

2:29 And the name of the wife of Abishur was Abihail, lii autem Semmei Nadab et Abisur 2:29 nomen uero uxo-

and she bare him Ahban, and Molid.

2:30 And the sons of Nadab; Seled, and Appaim: but ris Abisur Abiail quae peperit Ahobban et Molid 2:30 filii

Seled died without children.

autem Nadab fuerunt Saled et Apphaim mortuus est autem

2:31 And the sons of Appaim; Ishi. And the sons of Ishi; Saled absque liberis 2:31 filius uero Apphaim Iesi qui Iesi

Sheshan. And the children of Sheshan; Ahlai.

2:32 And the sons of Jada the brother of Shammai; Jether, genuit Sesan porro Sesan genuit Oholi 2:32 filii autem Iada

and Jonathan: and Jether died without children.

fratris Semmei Iether et Ionathan sed et Iether mortuus est

2:33 And the sons of Jonathan; Peleth, and Zaza. These absque liberis 2:33 porro Ionathan genuit Phaleth et Ziza

were the sons of Jerahmeel.

2:34 Now Sheshan had no sons, but daughters. And She- isti fuerunt filii Hieramehel 2:34 Sesan autem non habuit

shan had a servant, an Egyptian, whose name was Jarha.

2:35 And Sheshan gave his daughter to Jarha his servant filios sed filias et seruum aegyptium nomine Ieraa 2:35 de-

to wife; and she bare him Attai.

2:36 And Attai begat Nathan, and Nathan begat Zabad, ditque ei filiam suam uxorem quae peperit ei Eththei 2:36 2:37 And Zabad begat Ephlal, and Ephlal begat Obed, Eththei autem genuit Nathan et Nathan genuit Zabad 2:37 2:38 And Obed begat Jehu, and Jehu begat Azariah, Zabad quoque genuit Ophlal et Ophlal genuit Obed 2:38

2:39 And Azariah begat Helez, and Helez begat Eleasah, Obed genuit Ieu Ieu genuit Azariam 2:39 Azarias genuit

2:40 And Eleasah begat Sisamai, and Sisamai begat Shal- Helles Helles genuit Elasa 2:40 Elasa genuit Sisamoi Si-

lum,

2:41 And Shallum begat Jekamiah, and Jekamiah begat samoi genuit Sellum 2:41 Sellum genuit Icamian Icamian

Elishama.

2:42 Now the sons of Caleb the brother of Jerahmeel were, genuit Elisama 2:42 filii autem Chaleb fratris Hieramehel

Mesha his firstborn, which was the father of Ziph; and the sons of Mareshah the father of Hebron.

Mosa primogenitus eius ipse est pater Ziph et filii Maresa

2:43 And the sons of Hebron; Korah, and Tappuah, and patris Hebron 2:43 porro filii Hebron Core et Thapphu et

Rekem, and Shema. 2:44 And Shema begat Raham, the father of Jorkoam: and Recem et Samma 2:44 Samma autem genuit Raam patrem Rekem begat Shammai.

2:45 And the son of Shammai was Maon: and Maon was Iercaam et Recem genuit Semmei 2:45 filius Semmei Maon

the father of Bethzur.

2:46 And Ephah, Caleb’s concubine, bare Haran, and et Maon pater Bethsur 2:46 Epha autem concubina Cha-

Moza, and Gazez: and Haran begat Gazez.

leb peperit Arran et Musa et Gezez porro Arran genuit Ge-

2:47 And the sons of Jahdai; Regem, and Jotham, and Ge- zez 2:47 filii Iadai Regom et Iotham et Gesum et Phaleth sham, and Pelet, and Ephah, and Shaaph.

2:48 Maachah, Caleb’s concubine, bare Sheber, and et Epha et Saaph 2:48 concubina Chaleb Maacha peperit

Tirhanah.

2:49 She bare also Shaaph the father of Madmannah, Saber et Tharana 2:49 genuit autem Saaph pater Madmena

Sheva the father of Machbenah, and the father of Gibea: and the daughter of Caleb was Achsa.

Sue patrem Machbena et patrem Gabaa filia uero Chaleb fuit

2:50 These were the sons of Caleb the son of Hur, the Achsa 2:50 hii erant filii Chaleb filii Vr primogeniti Eph-

firstborn of Ephratah; Shobal the father of Kirjathjearim.

2:51 Salma the father of Bethlehem, Hareph the father of rata Sobal pater Cariathiarim 2:51 Salma pater Bethleem

Bethgader.

2:52 And Shobal the father of Kirjathjearim had sons; Ariph pater Bethgader 2:52 fuerunt autem filii Sobal pat-

Haroeh, and half of the Manahethites.

2:53 And the families of Kirjathjearim; the Ithrites, and ris Cariathiarim qui uidebat dimidium requietionum 2:53 et

the Puhites, and the Shumathites, and the Mishraites; of them came the Zareathites, and the Eshtaulites,

de cognatione Cariathiarim Iethrei et Apphutei et Semathei

2:54 The sons of Salma; Bethlehem, and the Ne- et Maserei ex his egressi sunt Saraitae et Esthaolitae 2:54

tophathites, Ataroth, the house of Joab, and half of the Manahethites, the Zorites.

filii Salma Bethleem et Netophathi coronae domus Ioab et

Liber I Paralipomenon

dimidium requietionis Sarai 2:55 cognationes quoque scri- 2:55 And the families of the scribes which dwelt at Jabez;

the Tirathites, the Shimeathites, and Suchathites. These are the Kenites that came of Hemath, the father of the

barum habitantium in Iabis canentes atque resonantes et in house of Rechab. tabernaculis commorantes hii sunt Cinei qui uenerunt de calore patris domus Rechab

3:1 Dauid uero hos habuit filios qui ei nati sunt in Hebron 3:1 Now these were the sons of David, which were

born unto him in Hebron; the firstborn Amnon, of Ahinoam the Jezreelitess; the second Daniel, of Abigail the

primogenitum Amnon ex Achinaam Iezrahelitide secundum Carmelitess: Danihel de Abigail Carmelitide 3:2 tertium Absalom filium 3:2 The third, Absalom the son of Maachah the daughter

of Talmai king of Geshur: the fourth, Adonijah the son of Haggith:

Maacha filiae Tholmei regis Gessur quartum Adoniam filium Aggith 3:3 quintum Saphatiam ex Abital sextum Ieth- 3:3 The fifth, Shephatiah of Abital: the sixth, Ithream by

Eglah his wife.

raam de Egla uxore sua 3:4 sex ergo nati sunt ei in Hebron 3:4 These six were born unto him in Hebron; and there he

reigned seven years and six months: and in Jerusalem he reigned thirty and three years.

ubi regnauit septem annis et sex mensibus triginta autem et tribus annis regnauit in Hierusalem 3:5 porro in Hierusa- 3:5 And these were born unto him in Jerusalem; Shimea,

and Shobab, and Nathan, and Solomon, four, of Bathshua the daughter of Ammiel:

lem nati sunt ei filii Samaa et Sobab et Nathan et Salomon quattuor de Bethsabee filia Amihel 3:6 Iebaar quoque et 3:6 Ibhar also, and Elishama, and Eliphelet, Elisama 3:7 et Eliphalet et Noge et Napheg et Iaphie 3:8 3:7 And Nogah, and Nepheg, and Japhia,

3:8 And Elishama, and Eliada, and Eliphelet, nine.

necnon Elisama et Heliade et Eliphalet nouem 3:9 omnes _{3:9 These were all the sons of David, beside the sons of}

the concubines, and Tamar their sister.

hii filii Dauid absque filiis concubinarum habuerunt sororem Thamar 3:10 filius autem Salomonis Roboam cuius Abia fi- 3:10 And Solomon’s son was Rehoboam, Abia his son,

Asa his son, Jehoshaphat his son,

lius genuit Asa de hoc quoque natus est Iosaphat 3:11 pater 3:11 Joram his son, Ahaziah his son, Joash his son, Ioram qui Ioram genuit Ohoziam ex quo ortus est Ioas 3:12 3:12 Amaziah his son, Azariah his son, Jotham his son, et huius Amasias filius genuit Azariam porro Azariae filius Ioatham 3:13 procreauit Achaz patrem Ezechiae de quo na- 3:13 Ahaz his son, Hezekiah his son, Manasseh his son, tus est Manasses 3:14 sed et Manasses genuit Amon patrem 3:14 Amon his son, Josiah his son. Iosiae 3:15 filii autem Iosiae fuerunt primogenitus Iohanan 3:15 And the sons of Josiah were, the firstborn Johanan,

the second Jehoiakim, the third Zedekiah, the fourth Shallum.

secundus Ioacim tertius Sedecias quartus Sellum 3:16 de

3:16 And the sons of Jehoiakim: Jeconiah his son, Zedekiah his son.

Ioacim natus est Iechonias et Sedecias 3:17 filii Iechoniae 3:17 And the sons of Jeconiah; Assir, Salathiel his son, fuerunt Asir Salathihel 3:18 Melchiram Phadaia Sennaser 3:18 Malchiram also, and Pedaiah, and Shenazar, Je-

camiah, Hoshama, and Nedabiah.

et Iecemia Sama et Nadabia 3:19 de Phadaia orti sunt Zo- 3:19 And the sons of Pedaiah were, Zerubbabel, and

Shimei: and the sons of Zerubbabel; Meshullam, and Hananiah, and Shelomith their sister:

robabel et Semei Zorobabel genuit Mosollam Ananiam et

Salomith sororem eorum 3:20 Asabamque et Ohol et Ba- 3:20 And Hashubah, and Ohel, and Berechiah, and Hasa-

diah, Jushabhesed, five.

rachiam et Asadiam Iosabesed quinque 3:21 filius autem 3:21 And the sons of Hananiah; Pelatiah, and Jesaiah: the sons of Rephaiah, the sons of Arnan, the sons of Obadiah, the sons of Shechaniah.

Ananiae Phaltias pater Ieseiae cuius filius Raphaia huius quoque filius Arnam de quo natus est Obdia cuius filius fuit

Biblia Sacra Vulgata

3:22 And the sons of Shechaniah; Shemaiah: and the sons Sechenia 3:22 filius Secheniae Semeia cuius filii Attus et

of Shemaiah; Hattush, and Igeal, and Bariah, and Neariah, and Shaphat, six.

Iegal et Baria et Naaria et Saphat sex numero 3:23 filius

3:23 And the sons of Neariah; Elioenai, and Hezekiah, and Azrikam, three.

3:24 And the sons of Elioenai were, Hodaiah, and Naariae Helioenai et Ezechias et Ezricam tres 3:24 filii He-

Eliashib, and Pelaiah, and Akkub, and Johanan, and Dalaiah, and Anani, seven.

lioenai Oduia et Heliasub et Pheleia et Accub et Iohanan et Dalaia et Anani septem

4:1 The sons of Judah; Pharez, Hezron, and Carmi, and 4:1 filii Iuda Phares Esrom et Carmi et Vr et Subal 4:2 Hur, and Shobal. 4:2 And Reaiah the son of Shobal begat Jahath; and Jahath begat Ahumai, and Lahad. These are the families of the Zorathites.

Reaia uero filius Subal genuit Ieth de quo nati sunt Ahimai

4:3 And these were of the father of Etam; Jezreel, and et Laed hae cognationes Sarathi 4:3 ista quoque stirps He-

Ishma, and Idbash: and the name of their sister was Hazelelponi:

tam Iezrahel et Iesema et Iedebos nomenque sororis eorum

4:4 And Penuel the father of Gedor, and Ezer the father Asalelphuni 4:4 Phunihel autem pater Gedor et Ezer pa-

of Hushah. These are the sons of Hur, the firstborn of Ephratah, the father of Bethlehem.

ter Osa isti sunt filii Vr primogeniti Ephrata patris Bethleem

4:5 And Ashur the father of Tekoa had two wives, Helah 4:5 Asur uero patris Thecue erant duae uxores Halaa et Na-

and Naarah.

4:6 And Naarah bare him Ahuzam, and Hepher, and ara 4:6 peperit autem ei Naara Oozam et Epher et Themani

Temeni, and Haahashtari. These were the sons of Naarah.

4:7 And the sons of Helah were, Zereth, and Jezoar, and et Asthari isti sunt filii Naara 4:7 porro filii Halaa Sereth

Ethnan. 4:8 And Coz begat Anub, and Zobebah, and the families Isaar et Ethnan 4:8 Cos autem genuit Anob et Sobaba et of Aharhel the son of Harum.

4:9 And Jabez was more honourable than his brethren: and cognationem Aral filii Arum 4:9 fuit autem Iabes inclitus

his mother called his name Jabez, saying, Because I bare him with sorrow.

prae fratribus suis et mater eius uocauit nomen illius Iabes

4:10 And Jabez called on the God of Israel, saying, Oh that dicens quia peperi eum in dolore 4:10 inuocauit uero Iabes

thou wouldest bless me indeed, and enlarge my coast, and that thine hand might be with me, and that thou wouldest keep me from evil, that it may not grieve me! And God granted him that which he requested.

Deum Israhel dicens si benedicens benedixeris mihi et dilataueris terminos meos et fuerit manus tua mecum et feceris me a malitia non opprimi et praestitit Deus quae precatus

4:11 And Chelub the brother of Shuah begat Mehir, which est 4:11 Chaleb autem frater Suaa genuit Machir qui fuit

was the father of Eshton.

4:12 And Eshton begat Bethrapha, and Paseah, and Tehin- pater Esthon 4:12 porro Esthon genuit Bethrapha et Phesse

nah the father of Irnahash. These are the men of Rechah.

4:13 And the sons of Kenaz; Othniel, and Seraiah: and the et Thena patrem urbis Naas hii sunt uiri Recha 4:13 filii au-

sons of Othniel; Hathath.

tem Cenez Othonihel et Saraia porro filii Othonihel Athath

4:14 And Meonothai begat Ophrah: and Seraiah begat 4:14 et Maonathi genuit Ophra Saraias autem genuit Ioab pat-

Joab, the father of the valley of Charashim; for they were craftsmen.

rem uallis Artificum ibi quippe artifices erant 4:15 filii uero

4:15 And the sons of Caleb the son of Jephunneh; Iru, Elah, and Naam: and the sons of Elah, even Kenaz.

Chaleb filii Iephonne Hir et Hela et Nahem filiique Hela et

4:16 And the sons of Jehaleleel; Ziph, and Ziphah, Tiria, Cenez 4:16 filii quoque Iallelel Ziph et Zipha Thiria et As-

and Asareel.

4:17 And the sons of Ezra were, Jether, and Mered, and rahel 4:17 et filii Ezra Iether et Mered et Epher et Ialon ge-

Epher, and Jalon: and she bare Miriam, and Shammai, and Ishbah the father of Eshtemoa.

nuitque Mariam et Sammai et Iesba patrem Esthamo 4:18

4:18 And his wife Jehudijah bare Jered the father of Gedor, and Heber the father of Socho, and Jekuthiel the father of Zanoah. And these are the sons of Bithiah the daughter of Pharaoh, which Mered took.

uxor quoque eius Iudaia peperit Iared patrem Gedor et Heber patrem Soccho et Hicuthihel patrem Zano hii autem filii

Liber I Paralipomenon

Beththiae filiae Pharaonis quam accepit Mered 4:19 et filii 4:19 And the sons of his wife Hodiah the sister of Na-

ham, the father of Keilah the Garmite, and Eshtemoa the Maachathite.

uxoris Odaiae sororis Naham patris Ceila Garmi et Esthamo qui fuit de Machathi 4:20 filii quoque Simon Amnon et Re- 4:20 And the sons of Shimon were, Amnon, and Rinnah,

Benhanan, and Tilon. And the sons of Ishi were, Zoheth, and Benzoheth.

na filius Anan et Thilon et filii Iesi Zoeth et Benzoeth 4:21

4:21 The sons of Shelah the son of Judah were, Er the father of Lecah, and Laadah the father of Mareshah, and the families of the house of them that wrought fine linen, of the house of Ashbea,

filii Sela filii Iuda Her pater Lecha et Laada pater Maresa et cognationes Domus operantium byssum in domo Iuramenti 4:22 et Qui stare fecit solem uirique Mendacii et Securus et 4:22 And Jokim, and the men of Chozeba, and Joash, and

Saraph, who had the dominion in Moab, and Jashubilehem. And these are ancient things.

Incendens qui principes fuerunt in Moab et qui reuersi sunt in Leem haec autem uerba uetera 4:23 hii sunt figuli habi- 4:23 These were the potters, and those that dwelt among

plants and hedges: there they dwelt with the king for his work.

tantes in plantationibus et in praesepibus apud regem in operibus eius commoratique sunt ibi 4:24 filii Symeon Namu- 4:24 The sons of Simeon were, Nemuel, and Jamin, Jarib,

Zerah, and Shaul:

hel et Iamin Iarib Zara Saul 4:25 Sellum filius eius Mabsam 4:25 Shallum his son, Mibsam his son, Mishma his son. filius eius Masma filius eius 4:26 filii Masma Amuhel filius 4:26 And the sons of Mishma; Hamuel his son, Zacchur

his son, Shimei his son.

eius Zacchur filius eius Semei filius eius 4:27 filii Semei 4:27 And Shimei had sixteen sons and six daughters: but

his brethren had not many children, neither did all their family multiply, like to the children of Judah.

sedecim et filiae sex fratres autem eius non habuerunt filios multos et uniuersa cognatio non potuit adaequare summam filiorum Iuda 4:28 habitauerunt autem in Bersabee et Mo- 4:28 And they dwelt at Beersheba, and Moladah, and Haz-

arshual,

lada et Asarsual 4:29 et in Ballaa et in Asom et in Tholad 4:29 And at Bilhah, and at Ezem, and at Tolad, 4:30 et in Bathuhel et in Orma et in Siceleg 4:31 et in Beth- 4:30 And at Bethuel, and at Hormah, and at Ziklag,

4:31 And at Bethmarcaboth, and Hazarsusim, and at Bethbirei, and at Shaaraim. These were their cities unto the reign of David.

marchaboth et in Asarsusim et in Bethberai et in Saarim hae ciuitates eorum usque ad regem Dauid 4:32 uillae quoque 4:32 And their villages were, Etam, and Ain, Rimmon,

and Tochen, and Ashan, five cities:

eorum Etham et Aen et Remmon et Thochen et Asan ciuitates quinque 4:33 et uniuersi uiculi eorum per circuitum 4:33 And all their villages that were round about the same

cities, unto Baal. These were their habitations, and their genealogy.

ciuitatum istarum usque ad Baal haec est habitatio eorum et sedum distributio 4:34 Masobab quoque et Iemlech et Iosa 4:34 And Meshobab, and Jamlech, and Joshah, the son of

Amaziah,

filius Amasiae 4:35 et Iohel et Ieu filius Iosabiae filii Sara- 4:35 And Joel, and Jehu the son of Josibiah, the son of

Seraiah, the son of Asiel,

iae filii Asihel 4:36 et Helioenai et Iacoba et Isuaia et Asaia 4:36 And Elioenai, and Jaakobah, and Jeshohaiah, and

Asaiah, and Adiel, and Jesimiel, and Benaiah,

et Adihel et Isimihel et Banaia 4:37 Ziza quoque filius Se- 4:37 And Ziza the son of Shiphi, the son of Allon, the son

of Jedaiah, the son of Shimri, the son of Shemaiah;

phei filii Allon filii Idaia filii Semri filii Samaia 4:38 isti 4:38 These mentioned by their names were princes in their

families: and the house of their fathers increased greatly.

sunt nominati principes in cognationibus suis et in domo adfinitatum suarum multiplicati sunt uehementer 4:39 et pro- 4:39 And they went to the entrance of Gedor, even unto

the east side of the valley, to seek pasture for their flocks.

fecti sunt ut ingrederentur in Gador usque ad orientem uallis et ut quaererent pascua gregibus suis 4:40 inueneruntque 4:40 And they found fat pasture and good, and the land

was wide, and quiet, and peaceable; for they of Ham had dwelt there of old.

Biblia Sacra Vulgata

pascuas uberes et ualde bonas et terram latissimam et quie-

4:41 And these written by name came in the days of tam et fertilem in qua ante habitauerunt de stirpe Ham 4:41

Hezekiah king of Judah, and smote their tents, and the habitations that were found there, and destroyed them utterly unto this day, and dwelt in their rooms: because there was pasture there for their flocks.

hii ergo uenerunt quos supra descripsimus nominatim in diebus Ezechiae regis Iuda et percusserunt tabernacula eorum et habitatores qui inuenti fuerant ibi et deleuerunt eos usque in praesentem diem habitaueruntque pro eis quoniam uber-

4:42 And some of them, even of the sons of Simeon, five rimas ibidem pascuas reppererunt 4:42 de filiis quoque Sy-

hundred men, went to mount Seir, having for their captains Pelatiah, and Neariah, and Rephaiah, and Uzziel, the sons

of Ishi. meon abierunt in montem Seir uiri quingenti habentes prin-

cipes Phaltiam et Nahariam et Raphaiam et Ozihel filios Iesi

4:43 And they smote the rest of the Amalekites that were 4:43 et percusserunt reliquias quae euadere potuerant Ama-

escaped, and dwelt there unto this day.

lechitarum et habitauerunt ibi pro eis usque ad diem hanc

5:1 Now the sons of Reuben the firstborn of Israel, (for he 5:1 filii quoque Ruben primogeniti Israhel ipse quippe fuit

was the firstborn; but forasmuch as he defiled his father’s bed, his birthright was given unto the sons of Joseph the son of Israel: and the genealogy is not to be reckoned after the birthright.

primogenitus eius sed cum uiolasset torum patris sui data sunt primogenita eius filiis Ioseph filii Israhel et non est ille

5:2 For Judah prevailed above his brethren, and of him reputatus in primogenitum 5:2 porro Iudas qui erat fortissi-

came the chief ruler; but the birthright was Joseph’s:)

mus inter fratres suos de stirpe eius principes germinati sunt

5:3 The sons, I say, of Reuben the firstborn of Israel were, primogenita autem reputata sunt Ioseph 5:3 filii ergo Ru-Hanoch, and Pallu, Hezron, and Carmi.

ben primogeniti Israhel Enoch et Phallu Esrom et Charmi

5:4 The sons of Joel; Shemaiah his son, Gog his son, 5:4 filii Iohel Samaia filius eius Gog filius eius Semei filius

Shimei his son,

5:5 Micah his son, Reaia his son, Baal his son, eius 5:5 Micha filius eius Reeia filius eius Baal filius eius 5:6 Beerah his son, whom Tilgathpilneser king of Assyria 5:6 Beera filius eius quem captiuum duxit Theglathphalnasar

carried away captive: he was prince of the Reubenites. 5:7 And his brethren by their families, when the genealogy rex Assyriorum et fuit princeps in tribu Ruben 5:7 fratres of their generations was reckoned, were the chief, Jeiel, and Zechariah,

autem eius et uniuersa cognatio quando numerabantur per

5:8 And Bela the son of Azaz, the son of Shema, the son of familias suas habuerunt principes Ieihel et Zacchariam 5:8

Joel, who dwelt in Aroer, even unto Nebo and Baalmeon:

porro Bala filius Azaz filii Samma filii Iohel ipse habitauit

5:9 And eastward he inhabited unto the entering in of the in Aroer usque ad Nebo et Beelmeon 5:9 contra orientawilderness from the river Euphrates: because their cattle were multiplied in the land of Gilead.

lem quoque plagam habitauit usque ad introitum heremi et flumen Eufraten multum quippe iumentorum numerum pos-

5:10 And in the days of Saul they made war with the Ha- sidebat in terra Galaad 5:10 in diebus autem Saul proeliati

garites, who fell by their hand: and they dwelt in their tents throughout all the east land of Gilead.

sunt contra Agareos et interfecerunt illos habitaueruntque pro eis in tabernaculis eorum in omni plaga quae respicit ad

5:11 And the children of Gad dwelt over against them, in orientem Galaad 5:11 filii uero Gad e regione eorum habi-

the land of Bashan unto Salcah:

5:12 Joel the chief, and Shapham the next, and Jaanai, and tauerunt in terra Basan usque Selcha 5:12 Iohel in capite et

Shaphat in Bashan.

Liber I Paralipomenon

Saphan secundus Ianai autem et Saphat in Basan 5:13 fra- 5:13 And their brethren of the house of their fathers

were, Michael, and Meshullam, and Sheba, and Jorai, and Jachan, and Zia, and Heber, seven.

tres uero eorum secundum domos cognationum suarum Michahel et Mosollam et Sebe et Iori et Iachan et Zie et Heber septem 5:14 hii filii Abiahil filii Vri filii Iaro filii Galaad fi- 5:14 These are the children of Abihail the son of Huri, the

son of Jaroah, the son of Gilead, the son of Michael, the son of Jeshishai, the son of Jahdo, the son of Buz;

lii Michahel filii Iesesi filii Ieddo filii Buz 5:15 fratres quo-

5:15 Ahi the son of Abdiel, the son of Guni, chief of the house of their fathers.

que filii Abdihel filii Guni princeps domus in familiis suis 5:16 et habitauerunt in Galaad et in Basan et in uiculis eius et 5:16 And they dwelt in Gilead in Bashan, and in her towns,

and in all the suburbs of Sharon, upon their borders.

in cunctis suburbanis Saron usque ad terminos 5:17 omnes 5:17 All these were reckoned by genealogies in the days

of Jotham king of Judah, and in the days of Jeroboam king of Israel.

hii numerati sunt in diebus Ioatham regis Iuda et in diebus Hieroboam regis Israhel 5:18 filii Ruben et Gad et dimi- 5:18 The sons of Reuben, and the Gadites, and half the

tribe of Manasseh, of valiant men, men able to bear buckler and sword, and to shoot with bow, and skilful in war, were four and forty thousand seven hundred and threescore, that went out to the war.

diae tribus Manasse uiri bellatores scuta portantes et gladios et tendentes arcum eruditique ad proelia quadraginta quattuor milia et septingenti sexaginta procedentes ad pugnam 5:19 dimicauerunt contra Agarenos Iturei uero et Naphei et 5:19 And they made war with the Hagarites, with Jetur,

and Nephish, and Nodab.

Nodab 5:20 praebuerunt eis auxilium traditique sunt in ma- 5:20 And they were helped against them, and the Hagarites

were delivered into their hand, and all that were with them: for they cried to God in the battle, and he was intreated of

nus eorum Agareni et uniuersi qui fuerant cum eis quia De- them; because they put their trust in him. um inuocauerunt cum proeliarentur et exaudiuit eos eo quod credidissent in eum 5:21 ceperuntque omnia quae possede- 5:21 And they took away their cattle; of their camels fifty

thousand, and of sheep two hundred and fifty thousand, and of asses two thousand, and of men an hundred thou-

rant camelorum quinquaginta milia et ouium ducenta quin- sand. quaginta milia asinos duo milia et animas hominum centum milia 5:22 uulnerati autem multi corruerunt fuit enim bel- 5:22 For there fell down many slain, because the war was

of God. And they dwelt in their steads until the captivity.

lum Domini habitaueruntque pro eis usque ad transmigrationem 5:23 filii quoque dimidiae tribus Manasse possederunt 5:23 And the children of the half tribe of Manasseh dwelt

in the land: they increased from Bashan unto Baalhermon and Senir, and unto mount Hermon.

terram a finibus Basan usque Baalhermon et Sanir et mon-

tem Hermon ingens quippe numerus erat 5:24 et hii fuerunt 5:24 And these were the heads of the house of their fathers, even Epher, and Ishi, and Eliel, and Azriel, and Jeremiah, and Hodaviah, and Jahdiel, mighty men of valour, famous

principes domus cognationis eorum Epher et Iesi et Heli- men, and heads of the house of their fathers. hel Ezrihel et Hieremia et Odoia et Iedihel uiri fortissimi et potentes et nominati duces in familiis suis 5:25 relique- 5:25 And they transgressed against the God of their fa-

thers, and went a whoring after the gods of the people of the land, whom God destroyed before them.

runt autem Deum patrum suorum et fornicati sunt post deos populorum terrae quos abstulit Dominus coram eis 5:26 et 5:26 And the God of Israel stirred up the spirit of Pul king

of Assyria, and the spirit of Tilgathpilneser king of Assyria, and he carried them away, even the Reubenites, and the Gadites, and the half tribe of Manasseh, and brought them unto Halah, and Habor, and Hara, and to the river Gozan, unto this day.

suscitauit Deus Israhel spiritum Ful regis Assyriorum et spiritum Theglathphalnasar regis Assur et transtulit Ruben et Gad et dimidium tribus Manasse et adduxit eos in Alae et

Abor et Ara et fluuium Gozan usque ad diem hanc

6:1 The sons of Levi; Gershon, Kohath, and Merari. 6:1 filii Leui Gersom Caath Merari 6:2 filii Caath Am-

6:2 And the sons of Kohath; Amram, Izhar, and Hebron, and Uzziel.

ram Isaar Hebron et Ozihel 6:3 filii Amram Aaron Moses

6:3 And the children of Amram; Aaron, and Moses, and Miriam. The sons also of Aaron; Nadab, and Abihu,

Eleazar, and Ithamar. et Maria filii Aaron Nadab et Abiu Eleazar et Ithamar 6:4

6:4 Eleazar begat Phinehas, Phinehas begat Abishua,

6:5 And Abishua begat Bukki, and Bukki begat Uzzi, Eleazar genuit Finees et Finees genuit Abisue 6:5 Abisue 6:6 And Uzzi begat Zerahiah, and Zerahiah begat uero genuit Bocci et Bocci genuit Ozi 6:6 Ozi genuit Zara-

Meraioth,

6:7 Meraioth begat Amariah, and Amariah begat Ahitub, iam et Zaraias genuit Meraioth 6:7 porro Meraioth genuit 6:8 And Ahitub begat Zadok, and Zadok begat Ahimaaz, Amariam et Amarias genuit Ahitob 6:8 Ahitob genuit Sa-6:9 And Ahimaaz begat Azariah, and Azariah begat Jo- doc Sadoc genuit Achimaas 6:9 Achimaas genuit Azariam

hanan,

6:10 And Johanan begat Azariah, (he it is that executed Azarias genuit Iohanan 6:10 Iohanan genuit Azariam ipse

the priest’s office in the temple that Solomon built in Jerusalem:)

est qui sacerdotio functus est in domo quam aedificauit Salo-

6:11 And Azariah begat Amariah, and Amariah begat Ahi- mon in Hierusalem 6:11 genuit autem Azarias Amariam et tub,

6:12 And Ahitub begat Zadok, and Zadok begat Shallum, Amarias genuit Ahitob 6:12 Ahitob genuit Sadoc et Sadoc 6:13 And Shallum begat Hilkiah, and Hilkiah begat genuit Sellum 6:13 Sellum genuit Helciam et Helcias ge-

Azariah,

6:14 And Azariah begat Seraiah, and Seraiah begat Je- nuit Azariam 6:14 Azarias genuit Saraiam et Saraias genuit

hozadak,

6:15 And Jehozadak went into captivity, when the LORD Iosedec 6:15 porro Iosedec egressus est quando transtulit

carried away Judah and Jerusalem by the hand of Nebuchadnezzar.

Dominus Iudam et Hierusalem per manus Nabuchodonosor

6:16 The sons of Levi; Gershom, Kohath, and Merari. 6:16 filii ergo Leui Gersom Caath et Merari 6:17 et haec

6:17 And these be the names of the sons of Gershom; Libni, and Shimei.

nomina filiorum Gersom Lobeni et Semei 6:18 filii Caath

6:18 And the sons of Kohath were, Amram, and Izhar, and Hebron, and Uzziel.

Amram et Isaar et Hebron et Ozihel 6:19 filii Merari Mo-

6:19 The sons of Merari; Mahli, and Mushi. And these are the families of the Levites according to their fathers.

oli et Musi hae autem cognationes Leui secundum familias

6:20 Of Gershom; Libni his son, Jahath his son, Zimmah eorum 6:20 Gersom Lobeni filius eius Iaath filius eius Za-

his son, 6:21 Joah his son, Iddo his son, Zerah his son, Jeaterai his mma filius eius 6:21 Ioaa filius eius Addo filius eius Zara son.

6:22 The sons of Kohath; Amminadab his son, Korah his filius eius Iethrai filius eius 6:22 filii Caath Aminadab filius

son, Assir his son,

6:23 Elkanah his son, and Ebiasaph his son, and Assir his eius Core filius eius Asir filius eius 6:23 Helcana filius eius

son,

6:24 Tahath his son, Uriel his son, Uzziah his son, and Abiasaph filius eius Asir filius eius 6:24 Thaath filius eius

Shaul his son.

6:25 And the sons of Elkanah; Amasai, and Ahimoth. Vrihel filius eius Ozias filius eius Saul filius eius 6:25 fi-

6:26 As for Elkanah: the sons of Elkanah; Zophai his son, lii Helcana Amasai et Ahimoth 6:26 Helcana filii Helcana and Nahath his son,

6:27 Eliab his son, Jeroham his son, Elkanah his son. Sophai filius eius Naath filius eius 6:27 Heliab filius eius 6:28 And the sons of Samuel; the firstborn Vashni, and Hieroam filius eius Helcana filius eius 6:28 filii Samuhel

Abiah. 6:29 The sons of Merari; Mahli, Libni his son, Shimei his primogenitus Vasseni et Abia 6:29 filii autem Merari Moson, Uzza his son,

6:30 Shimea his son, Haggiah his son, Asaiah his son. oli Lobeni filius eius Semei filius eius Oza filius eius 6:30 6:31 And these are they whom David set over the service Samaa filius eius Aggia filius eius Asaia filius eius 6:31 isti

of song in the house of the LORD, after that the ark had rest.

Liber I Paralipomenon

sunt quos constituit Dauid super cantores domus Domini ex quo conlocata est arca 6:32 et ministrabant coram taberna- 6:32 And they ministered before the dwelling place of the

tabernacle of the congregation with singing, until Solomon had built the house of the LORD in Jerusalem: and then

culo testimonii canentes donec aedificaret Salomon domum they waited on their office according to their order. Domini in Hierusalem stabant autem iuxta ordinem suum in ministerio 6:33 hii uero sunt qui adsistebant cum filiis suis 6:33 And these are they that waited with their children.

Of the sons of the Kohathites: Heman a singer, the son of Joel, the son of Shemuel,

de filiis Caath Heman cantor filius Iohel filii Samuhel 6:34

6:34 The son of Elkanah, the son of Jeroham, the son of Eliel, the son of Toah,

filii Helcana filii Hieroam filii Helihel filii Thou 6:35 filii_{6:35 The son of Zuph, the son of Elkanah, the son of Ma-}

hath, the son of Amasai,

Suph filii Helcana filii Maath filii Amasai 6:36 filii Helcana 6:36 The son of Elkanah, the son of Joel, the son of

Azariah, the son of Zephaniah,

filii Iohel filii Azariae filii Sophoniae 6:37 filii Thaath filii 6:37 The son of Tahath, the son of Assir, the son of Ebias-

aph, the son of Korah,

Asir filii Abiasaph filii Core 6:38 filii Isaar filii Caath fi- 6:38 The son of Izhar, the son of Kohath, the son of Levi,

the son of Israel.

lii Leui filii Israhel 6:39 et fratres eius Asaph qui stabat a 6:39 And his brother Asaph, who stood on his right hand,

even Asaph the son of Berachiah, the son of Shimea,

dextris eius Asaph filius Barachiae filii Samaa 6:40 filii Mi- 6:40 The son of Michael, the son of Baaseiah, the son of

Malchiah,

chahel filii Basiae filii Melchiae 6:41 filii Athnai filii Zara 6:41 The son of Ethni, the son of Zerah, the son of Adaiah, filii Adaia 6:42 filii Ethan filii Zamma filii Semei 6:43 fi- 6:42 The son of Ethan, the son of Zimmah, the son of

Shimei,

6:43 The son of Jahath, the son of Gershom, the son of Levi.

lii Ieth filii Gersom filii Leui 6:44 filii autem Merari frat-

6:44 And their brethren the sons of Merari stood on the left hand: Ethan the son of Kishi, the son of Abdi, the son of Malluch,

res eorum ad sinistram Ethan filius Cusi filii Abdi filii Maloch 6:45 filii Asabiae filii Amasiae filii Helciae 6:46 filii 6:45 The son of Hashabiah, the son of Amaziah, the son

of Hilkiah,

6:46 The son of Amzi, the son of Bani, the son of Shamer,

Amasai filii Bonni filii Somer 6:47 filii Mooli filii Musi fi-

6:47 The son of Mahli, the son of Mushi, the son of Merari, the son of Levi.

lii Merari filii Leui 6:48 fratres quoque eorum Leuitae qui

6:48 Their brethren also the Levites were appointed unto all manner of service of the tabernacle of the house of God.

ordinati sunt in cunctum ministerium tabernaculi domus Domini 6:49 Aaron uero et filii eius adolebant incensum super 6:49 But Aaron and his sons offered upon the altar of the

burnt offering, and on the altar of incense, and were appointed for all the work of the place most holy, and to make an atonement for Israel, according to all that Moses the servant of God had commanded.

altare holocausti et super altare thymiamatis in omne opus sancti sanctorum et ut precarentur pro Israhel iuxta omnia quae praecepit Moses seruus Dei 6:50 hii sunt autem fi- 6:50 And these are the sons of Aaron; Eleazar his son,

Phinehas his son, Abishua his son,

lii Aaron Eleazar filius eius Finees filius eius Abisue filius eius 6:51 Bocci filius eius Ozi filius eius Zaraia filius eius 6:51 Bukki his son, Uzzi his son, Zerahiah his son, 6:52 Meraioth filius eius Amaria filius eius Ahitob filius eius 6:52 Meraioth his son, Amariah his son, Ahitub his son, 6:53 Sadoc filius eius Achimaas filius eius 6:54 et haec ha- 6:53 Zadok his son, Ahimaaz his son.

6:54 Now these are their dwelling places throughout their castles in their coasts, of the sons of Aaron, of the families of the Kohathites: for theirs was the lot.

bitacula eorum per uicos atque confinia filiorum scilicet Aaron iuxta cognationes Caathitarum ipsis enim sorte contigerat 6:55 dederunt igitur eis Hebron in terra Iuda et subur- 6:55 And they gave them Hebron in the land of Judah, and

the suburbs thereof round about it.

bana eius per circuitum 6:56 agros autem ciuitatis et uil- 6:56 But the fields of the city, and the villages thereof, they

gave to Caleb the son of Jephunneh.

las Chaleb filio Iephonne 6:57 porro filiis Aaron dederunt 6:57 And to the sons of Aaron they gave the cities of Ju-

dah, namely, Hebron, the city of refuge, and Libnah with her suburbs, and Jattir, and Eshtemoa, with their suburbs,

Biblia Sacra Vulgata

ciuitates ad confugiendum Hebron et Lobna et suburbana

6:58 And Hilen with her suburbs, Debir with her suburbs, eius 6:58 Iether quoque et Esthmo cum suburbanis suis sed

6:59 And Ashan with her suburbs, and Bethshemesh with et Helon et Dabir cum suburbanis suis 6:59 Asan quoque her suburbs: 6:60 And out of the tribe of Benjamin; Geba with her sub- et Bethsemes et suburbana eorum 6:60 de tribu autem Beurbs, and Alemeth with her suburbs, and Anathoth with her suburbs. All their cities throughout their families were

thirteen cities. niamin Gabee et suburbana eius et Almath cum suburbanis

suis Anathoth quoque cum suburbanis suis omnes ciuitates

6:61 And unto the sons of Kohath, which were left of the tredecim per cognationes suas 6:61 filiis autem Caath resi-

family of that tribe, were cities given out of the half tribe, namely, out of the half tribe of Manasseh, by lot, ten cities.

duis de cognatione sua dederunt ex dimidia tribu Manasse

6:62 And to the sons of Gershom throughout their families in possessionem urbes decem 6:62 porro filiis Gersom per out of the tribe of Issachar, and out of the tribe of Asher, and out of the tribe of Naphtali, and out of the tribe of

Manasseh in Bashan, thirteen cities. cognationes suas de tribu Isachar et de tribu Aser et de tribu

6:63 Unto the sons of Merari were given by lot, throughout Nepthali et de tribu Manasse in Basan urbes tredecim 6:63 their families, out of the tribe of Reuben, and out of the tribe of Gad, and out of the tribe of Zebulun, twelve cities.

filiis autem Merari per cognationes suas de tribu Ruben et de tribu Gad et de tribu Zabulon dederunt sorte ciuitates duode-

6:64 And the children of Israel gave to the Levites these cim 6:64 dederunt quoque filii Israhel Leuitis ciuitates et

cities with their suburbs. 6:65 And they gave by lot out of the tribe of the children of suburbana earum 6:65 dederuntque per sortem ex tribu fi-Judah, and out of the tribe of the children of Simeon, and out of the tribe of the children of Benjamin, these cities,

which are called by their names. liorum Iuda et ex tribu filiorum Symeon et ex tribu filiorum

6:66 And the residue of the families of the sons of Kohath Beniamin urbes has quas uocauerunt nominibus suis 6:66 had cities of their coasts out of the tribe of Ephraim.

et his qui erant ex cognatione filiorum Caath fueruntque ci-

6:67 And they gave unto them, of the cities of refuge, uitates in terminis eorum de tribu Ephraim 6:67 dederunt

Shechem in mount Ephraim with her suburbs; they gave also Gezer with her suburbs,

ergo eis urbes ad confugiendum Sychem cum suburbanis

6:68 And Jokmeam with her suburbs, and Bethhoron with suis in monte Ephraim et Gazer cum suburbanis suis 6:68

her suburbs,

Hicmaam quoque cum suburbanis suis et Bethoron similiter

6:69 And Aijalon with her suburbs, and Gathrimmon with 6:69 necnon et Helon cum suburbanis suis et Gethremmon in

her suburbs: 6:70 And out of the half tribe of Manasseh; Aner with her eundem modum 6:70 porro ex dimidia tribu Manasse Aner suburbs, and Bileam with her suburbs, for the family of the remnant of the sons of Kohath.

et suburbana eius Balaam et suburbana eius his uidelicet qui

6:71 Unto the sons of Gershom were given out of the fam- de cognatione filiorum Caath reliqui erant 6:71 filiis autem

ily of the half tribe of Manasseh, Golan in Bashan with her suburbs, and Ashtaroth with her suburbs:

Gersom de cognatione dimidiae tribus Manasse Gaulon in Basan et suburbana eius et Astharoth cum suburbanis suis

6:72 And out of the tribe of Issachar; Kedesh with her 6:72 de tribu Isachar Cedes et suburbana eius et Dabereth

suburbs, Daberath with her suburbs,

6:73 And Ramoth with her suburbs, and Anem with her cum suburbanis suis 6:73 Ramoth quoque et suburbana il-

suburbs:

6:74 And out of the tribe of Asher; Mashal with her sub- lius et Anem cum suburbanis suis 6:74 de tribu uero Aser

urbs, and Abdon with her suburbs,

6:75 And Hukok with her suburbs, and Rehob with her Masal cum suburbanis suis et Abdon similiter 6:75 Acac

suburbs:

6:76 And out of the tribe of Naphtali; Kedesh in Galilee quoque et suburbana eius et Roob cum suburbanis suis 6:76

with her suburbs, and Hammon with her suburbs, and Kirjathaim with her suburbs.

Liber I Paralipomenon

porro de tribu Nepthali Cedes in Galilea et suburbana eius Amon cum suburbanis suis et Cariathaim et suburbana eius 6:77 filiis autem Merari residuis de tribu Zabulon Remmono 6:77 Unto the rest of the children of Merari were given out

of the tribe of Zebulun, Rimmon with her suburbs, Tabor with her suburbs:

et suburbana eius et Thabor cum suburbanis suis 6:78 trans

6:78 And on the other side Jordan by Jericho, on the east side of Jordan, were given them out of the tribe of Reuben, Bezer in the wilderness with her suburbs, and Jahzah with her suburbs,

Iordanem quoque ex aduerso Hiericho contra orientem Iordanis de tribu Ruben Bosor in solitudine cum suburbanis suis et Iasa cum suburbanis suis 6:79 Cademoth quoque et su- 6:79 Kedemoth also with her suburbs, and Mephaath with

her suburbs:

burbana eius et Miphaath cum suburbanis suis 6:80 necnon 6:80 And out of the tribe of Gad; Ramoth in Gilead with

her suburbs, and Mahanaim with her suburbs,

de tribu Gad Ramoth in Galaad et suburbana eius et Manaim cum suburbanis suis 6:81 sed et Esbon cum suburbanis eius 6:81 And Heshbon with her suburbs, and Jazer with her

suburbs.

et Iezer cum suburbanis suis

7:1 porro filii Isachar Thola et Phua Iasub et Samaron 7:1 Now the sons of Issachar were, Tola, and Puah, Jashub,

and Shimrom, four.

quattuor 7:2 filii Thola Ozi et Raphaia et Ierihel et Iemai et 7:2 And the sons of Tola; Uzzi, and Rephaiah, and Jeriel,

and Jahmai, and Jibsam, and Shemuel, heads of their father’s house, to wit, of Tola: they were valiant men of might in their generations; whose number was in the days of David two and twenty thousand and six hundred.

Iebsem et Samuhel principes per domos cognationum suarum de stirpe Thola uiri fortissimi numerati sunt in diebus Dauid uiginti duo milia sescenti 7:3 filii Ozi Iezraia de quo 7:3 And the sons of Uzzi; Izrahiah: and the sons of

Izrahiah; Michael, and Obadiah, and Joel, Ishiah, five: all of them chief men.

nati sunt Michahel et Obadia et Iohel et Iesia quinque om-

nes principes 7:4 cumque eis per familias et populos suos 7:4 And with them, by their generations, after the house of their fathers, were bands of soldiers for war, six and thirty thousand men: for they had many wives and sons.

accincti ad proelium uiri fortissimi triginta sex milia multas enim habuere uxores et filios 7:5 fratresque eorum per om- 7:5 And their brethren among all the families of Issachar

were valiant men of might, reckoned in all by their genealogies fourscore and seven thousand.

nem cognationem Isachar robustissimi ad pugnandum octoginta septem milia numerati sunt 7:6 Beniamin Bale et Bo- 7:6 The sons of Benjamin; Bela, and Becher, and Jediael,

three.

chor et Iadihel tres 7:7 filii Bale Esbon et Ozi et Ozihel et 7:7 And the sons of Bela; Ezbon, and Uzzi, and Uzziel,

and Jerimoth, and Iri, five; heads of the house of their fathers, mighty men of valour; and were reckoned by their

Ierimoth et Vrai quinque principes familiarum et ad pugnan- genealogies twenty and two thousand and thirty and four. dum robustissimi numerus autem eorum uiginti duo milia et triginta quattuor 7:8 porro filii Bochor Zamira et Ioas et 7:8 And the sons of Becher; Zemira, and Joash, and

Eliezer, and Elioenai, and Omri, and Jerimoth, and Abiah, and Anathoth, and Alameth. All these are the sons of

Eliezer et Helioenai et Amri et Ierimoth et Abia et Anathoth Becher. et Almathan omnes hii filii Bochor 7:9 numerati sunt autem 7:9 And the number of them, after their genealogy by their

generations, heads of the house of their fathers, mighty men of valour, was twenty thousand and two hundred.

per familias suas principes cognationum ad bella fortissimi uiginti milia et ducenti 7:10 porro filii Iadihel Balan filii au- 7:10 The sons also of Jediael; Bilhan: and the sons of Bil-

han; Jeush, and Benjamin, and Ehud, and Chenaanah, and Zethan, and Tharshish, and Ahishahar.

tem Balan Hieus et Beniamin et Ahoth et Chanana et Iothan et Tharsis et Haisaar 7:11 omnes hii filii Iadihel principes 7:11 All these the sons of Jediael, by the heads of their

fathers, mighty men of valour, were seventeen thousand and two hundred soldiers, fit to go out for war and battle.

cognationum suarum uiri fortissimi decem et septem milia

Biblia Sacra Vulgata

7:12 Shuppim also, and Huppim, the children of Ir, and et ducenti ad proelium procedentes 7:12 Sephan quoque et

Hushim, the sons of Aher. 7:13 The sons of Naphtali; Jahziel, and Guni, and Jezer, Apham filii Hir et Asim filii Aer 7:13 filii autem Nepthali and Shallum, the sons of Bilhah.

7:14 The sons of Manasseh; Ashriel, whom she bare: (but Iasihel et Guni et Asar et Sellum filii Balaa 7:14 porro fi-

his concubine the Aramitess bare Machir the father of Gilead:

lius Manasse Esrihel concubinaque eius syra peperit Machir

7:15 And Machir took to wife the sister of Huppim and patrem Galaad 7:15 Machir autem accepit uxores filiis suis

Shuppim, whose sister’s name was Maachah;) and the name of the second was Zelophehad: and Zelophehad had

daughters. Happhim et Sepham et habuit sororem nomine Maacha no-

men autem secundi Salphaad nataeque sunt Salphaad filiae

7:16 And Maachah the wife of Machir bare a son, and she 7:16 et peperit Maacha uxor Machir filium uocauitque no-

called his name Peresh; and the name of his brother was Sheresh; and his sons were Ulam and Rakem.

men eius Phares porro nomen fratris eius Sares et filii eius

7:17 And the sons of Ulam; Bedan. These were the sons Vlam et Recem 7:17 filius autem Vlam Badan hii sunt filii

of Gilead, the son of Machir, the son of Manasseh. 7:18 And his sister Hammoleketh bare Ishod, and Abiezer, Galaad filii Machir filii Manasse 7:18 soror autem eius Reand Mahalah.

7:19 And the sons of Shemidah were, Ahian, and gina peperit uirum Decorum et Abiezer et Moola 7:19 erant

Shechem, and Likhi, and Aniam. 7:20 And the sons of Ephraim; Shuthelah, and Bered his autem filii Semida Ahin et Sechem et Leci et Aniam 7:20 son, and Tahath his son, and Eladah his son, and Tahath his son,

filii autem Ephraim Suthala Bareth filius eius Thaath filius eius Elada filius eius Thaath filius eius et huius filius Za-

7:21 And Zabad his son, and Shuthelah his son, and Ezer, bad 7:21 et huius filius Suthala et huius filius Ezer et Elad

and Elead, whom the men of Gath that were born in that land slew, because they came down to take away their cat-

tle. occiderunt autem eos uiri Geth indigenae quia descenderant 7:22 And Ephraim their father mourned many days, and ut inuaderent possessiones eorum 7:22 luxit igitur Ephraim

his brethren came to comfort him.

pater eorum multis diebus et uenerunt fratres eius ut conso-

7:23 And when he went in to his wife, she conceived, and larentur eum 7:23 ingressusque est ad uxorem suam quae

bare a son, and he called his name Beriah, because it went evil with his house.

concepit et peperit filium et uocauit nomen eius Beria eo

7:24 (And his daughter was Sherah, who built Bethhoron quod in malis domus eius ortus esset 7:24 filia autem eius

the nether, and the upper, and Uzzensherah.)

fuit Sara quae aedificauit Bethoron inferiorem et superio-

7:25 And Rephah was his son, also Resheph, and Telah his rem et Ozensara 7:25 porro filius eius Rapha et Reseph et son, and Tahan his son.

7:26 Laadan his son, Ammihud his son, Elishama his son. Thale de quo natus est Thaan 7:26 qui genuit Laadan huius 7:27 Non his son, Jehoshuah his son. quoque filius Ammiud genuit Elisama 7:27 de quo ortus est

7:28 And their possessions and habitations were, Bethel Nun qui habuit filium Iosue 7:28 possessio autem eorum et and the towns thereof, and eastward Naaran, and westward Gezer, with the towns thereof; Shechem also and the towns

thereof, unto Gaza and the towns thereof: habitatio Bethel cum filiabus suis et contra orientem Noran

ad occidentalem plagam Gazer et filiae eius Sychem quoque

7:29 And by the borders of the children of Man- cum filiabus suis usque Aza et filias eius 7:29 iuxta filios

asseh, Bethshean and her towns, Taanach and her towns, Megiddo and her towns, Dor and her towns. In these dwelt

the children of Joseph the son of Israel. quoque Manasse Bethsan et filias eius Thanach et filias eius

Mageddo et filias eius Dor et filias eius in his habitauerunt

7:30 The sons of Asher; Imnah, and Isuah, and Ishuai, and filii Ioseph filii Israhel 7:30 filii Aser Iomna et Iesua et Isui

Beriah, and Serah their sister.

Liber I Paralipomenon

et Baria et Sara soror eorum 7:31 filii autem Baria Heber et 7:31 And the sons of Beriah; Heber, and Malchiel, who is

the father of Birzavith.

Melchihel ipse est pater Barzaith 7:32 Heber autem genuit 7:32 And Heber begat Japhlet, and Shomer, and Hotham,

and Shua their sister.

Iephlat et Somer et Otham et Suaa sororem eorum 7:33 filii 7:33 And the sons of Japhlet; Pasach, and Bimhal, and

Ashvath. These are the children of Japhlet.

Iephlat Phosech et Chamaal et Asoth hii filii Iephlat 7:34 7:34 And the sons of Shamer; Ahi, and Rohgah, Jehubbah,

and Aram.

porro filii Somer Ahi et Roaga et Iaba et Aram 7:35 filii 7:35 And the sons of his brother Helem; Zophah, and

Imna, and Shelesh, and Amal.

autem Helem fratris eius Supha et Iemna et Selles et Amal 7:36 filii Supha Sue Arnaphed et Sual et Beri et Iamra 7:37 7:36 The sons of Zophah; Suah, and Harnepher, and

Shual, and Beri, and Imrah,

7:37 Bezer, and Hod, and Shamma, and Shilshah, and Ithran, and Beera.

Bosor et Od et Samma et Salusa et Iethran et Bera 7:38 fi-

7:38 And the sons of Jether; Jephunneh, and Pispah, and

lii Iether Iephonne et Phaspha et Ara 7:39 filii autem Olla _{Ara.}

7:39 And the sons of Ulla; Arah, and Haniel, and Rezia.

Aree et Anihel et Resia 7:40 omnes hii filii Aser principes 7:40 All these were the children of Asher, heads of their

father’s house, choice and mighty men of valour, chief of the princes. And the number throughout the genealogy of them that were apt to the war and to battle was twenty and six thousand men.

cognationum electi atque fortissimi duces ducum numerus autem eorum aetatis quae apta esset ad bellum uiginti sex milia

8:1 Beniamin autem genuit Bale primogenitum suum As- 8:1 Now Benjamin begat Bela his firstborn, Ashbel the

second, and Aharah the third,

bal secundum Ohora tertium 8:2 Nuaha quartum et Rapha 8:2 Nohah the fourth, and Rapha the fifth. quintum 8:3 fueruntque filii Bale Addaor et Gera et Abiud 8:3 And the sons of Bela were, Addar, and Gera, and Abi-

hud,

8:4 Abisue quoque et Neman et Ahoe 8:5 sed et Gera et 8:4 And Abishua, and Naaman, and Ahoah,

8:5 And Gera, and Shephuphan, and Huram.

Sephuphan et Vram 8:6 hii sunt filii Aod principes cogna- _{8:6 And these are the sons of Ehud: these are the heads of}

the fathers of the inhabitants of Geba, and they removed them to Manahath:

tionum habitantium in Gabaa qui translati sunt in Manath 8:7 Nooman autem et Achia et Gera ipse transtulit eos et ge- 8:7 And Naaman, and Ahiah, and Gera, he removed them,

and begat Uzza, and Ahihud.

nuit Oza et Ahiud 8:8 porro Saarim genuit in regione Moab 8:8 And Shaharaim begat children in the country of Moab,

after he had sent them away; Hushim and Baara were his wives.

postquam dimisit Vsim et Bara uxores suas 8:9 genuit au-

8:9 And he begat of Hodesh his wife, Jobab, and Zibia, and Mesha, and Malcham,

tem de Edes uxore sua Iobab et Sebia et Mosa et Molchom 8:10 Iehus quoque et Sechia et Marma hii sunt filii eius prin- 8:10 And Jeuz, and Shachia, and Mirma. These were his

sons, heads of the fathers.

cipes in familiis suis 8:11 Meusim uero genuit Abitob et 8:11 And of Hushim he begat Abitub, and Elpaal. Elphaal 8:12 porro filii Elphaal Heber et Misaam et Samad 8:12 The sons of Elpaal; Eber, and Misham, and Shamed,

who built Ono, and Lod, with the towns thereof:

hic aedificauit Ono et Lod et filias eius 8:13 Bara autem et 8:13 Beriah also, and Shema, who were heads of the fa-

thers of the inhabitants of Aijalon, who drove away the inhabitants of Gath:

Samma principes cognationum habitantium in Aialon hii fugauerunt habitatores Geth 8:14 et Haio et Sesac et Ierimoth 8:14 And Ahio, Shashak, and Jeremoth, 8:15 et Zabadia et Arod et Eder 8:16 Michahel quoque et 8:15 And Zebadiah, and Arad, and Ader,

8:16 And Michael, and Ispah, and Joha, the sons of Beriah;

Iespha et Ioaa filii Baria 8:17 et Zabadia et Mosollam et

8:17 And Zebadiah, and Meshullam, and Hezeki, and Heber,

Ezeci et Heber 8:18 et Iesamari et Iezlia et Iobab filii El-

8:18 Ishmerai also, and Jezliah, and Jobab, the sons of Elpaal;

phaal 8:19 et Iacim et Zechri et Zabdi 8:20 et Helioenai et _{8:19 And Jakim, and Zichri, and Zabdi,}

8:20 And Elienai, and Zilthai, and Eliel,

Biblia Sacra Vulgata

8:21 And Adaiah, and Beraiah, and Shimrath, the sons of Selethai et Helihel 8:21 et Adaia et Baraia et Samarath filii

Shimhi;

8:22 And Ishpan, and Heber, and Eliel, Semei 8:22 et Iesphan et Heber et Helihel 8:23 et Abdon

8:23 And Abdon, and Zichri, and Hanan,

8:24 And Hananiah, and Elam, and Antothijah, et Zechri et Hanan 8:24 et Anania et Ailam et Anathothia 8:25 And Iphedeiah, and Penuel, the sons of Shashak; 8:25 et Iephdaia et Phanuhel filii Sesac 8:26 et Samsari et

8:26 And Shamsherai, and Shehariah, and Athaliah, 8:27 And Jaresiah, and Eliah, and Zichri, the sons of Jero- Sooria et Otholia 8:27 et Iersia et Helia et Zechri filii Ieham. 8:28 These were heads of the fathers, by their generations, roam 8:28 hii patriarchae et cognationum principes qui hachief men. These dwelt in Jerusalem.

8:29 And at Gibeon dwelt the father of Gibeon; whose bitauerunt in Hierusalem 8:29 in Gabaon autem habitaue-

wife’s name was Maachah:

8:30 And his firstborn son Abdon, and Zur, and Kish, and runt Abigabaon et nomen uxoris eius Maacha 8:30 filius-

Baal, and Nadab,

que eius primogenitus Abdon et Sur et Cis et Baal et Nadab

8:31 And Gedor, and Ahio, and Zacher. 8:31 Gedor quoque et Ahio et Zacher 8:32 et Macelloth ge-

8:32 And Mikloth begat Shimeah. And these also dwelt with their brethren in Jerusalem, over against them.

nuit Samaa habitaueruntque ex aduerso fratrum suorum in

8:33 And Ner begat Kish, and Kish begat Saul, and Saul Hierusalem cum fratribus suis 8:33 Ner autem genuit Cis

begat Jonathan, and Malchishua, and Abinadab, and Eshbaal.

et Cis genuit Saul porro Saul genuit Ionathan et Melchisuae

8:34 And the son of Jonathan was Meribbaal; and Merib- et Abinadab et Esbaal 8:34 filius autem Ionathan Merib-

baal begat Micah.

8:35 And the sons of Micah were, Pithon, and Melech, and baal et Meribbaal genuit Micha 8:35 filii Micha Phithon et

Tarea, and Ahaz.

8:36 And Ahaz begat Jehoadah; and Jehoadah begat Melech et Thara et Ahaz 8:36 et Ahaz genuit Ioada et Io-

Alemeth, and Azmaveth, and Zimri; and Zimri begat Moza,

ada genuit Almoth et Azmoth et Zamari porro Zamari genuit

8:37 And Moza begat Binea: Rapha was his son, Eleasah Mosa 8:37 et Mosa genuit Baana cuius filius fuit Rapha de

his son, Azel his son:

8:38 And Azel had six sons, whose names are these, quo ortus est Elasa qui genuit Asel 8:38 porro Asel sex filii

Azrikam, Bocheru, and Ishmael, and Sheariah, and Obadiah, and Hanan. All these were the sons of Azel.

fuere his nominibus Ezricam Bochru Ismahel Saria Abadia

8:39 And the sons of Eshek his brother were, Ulam his Anan omnes hii filii Asel 8:39 filii autem Esec fratris eius

firstborn, Jehush the second, and Eliphelet the third.

8:40 And the sons of Ulam were mighty men of valour, Vlam primogenitus et Vs secundus et Eliphalet tertius 8:40

archers, and had many sons, and sons’ sons, an hundred and fifty. All these are of the sons of Benjamin.

fueruntque filii Vlam uiri robustissimi et magno robore tendentes arcum et multos habentes filios ac nepotes usque ad centum quinquaginta omnes hii filii Beniamin

9:1 So all Israel were reckoned by genealogies; and, be- 9:1 uniuersus ergo Israhel dinumeratus est et summa eohold, they were written in the book of the kings of Israel and Judah, who were carried away to Babylon for their

transgression. rum scripta est in libro regum Israhel et Iuda translatique

9:2 Now the first inhabitants that dwelt in their possessions sunt in Babylonem propter delictum suum 9:2 qui autem in their cities were, the Israelites, the priests, Levites, and the Nethinims.

habitauerunt primi in possessionibus et in urbibus suis Isra-

9:3 And in Jerusalem dwelt of the children of Judah, and of hel et sacerdotes Leuitae et Nathinnei 9:3 commorati sunt

the children of Benjamin, and of the children of Ephraim, and Manasseh;

in Hierusalem de filiis Iuda et de filiis Beniamin de filiis

9:4 Uthai the son of Ammihud, the son of Omri, the son quoque Ephraim et Manasse 9:4 Othei filius Amiud filius

of Imri, the son of Bani, of the children of Pharez the son of Judah.

Emri filii Omrai filii Bonni de filiis Phares filii Iuda 9:5

9:5 And of the Shilonites; Asaiah the firstborn, and his sons.

Liber I Paralipomenon

et de Siloni Asaia primogenitus et filii eius 9:6 de filiis 9:6 And of the sons of Zerah; Jeuel, and their brethren, six

hundred and ninety.

autem Zara Ieuhel et fratres eorum sescenti nonaginta 9:7 9:7 And of the sons of Benjamin; Sallu the son of Meshul-

lam, the son of Hodaviah, the son of Hasenuah,

porro de filiis Beniamin Salo filius Mosollam filii Oduia filii Asana 9:8 et Iobania filius Hieroam et Hela filius Ozi 9:8 And Ibneiah the son of Jeroham, and Elah the son of

Uzzi, the son of Michri, and Meshullam the son of Shephathiah, the son of Reuel, the son of Ibnijah;

filii Mochori et Mosollam filius Saphatiae filii Rahuhel fi-

lii Iebaniae 9:9 et fratres eorum per familias suas nongenti 9:9 And their brethren, according to their generations, nine hundred and fifty and six. All these men were chief of the fathers in the house of their fathers.

quinquaginta sex omnes hii principes cognationum per do-

mos patrum suorum 9:10 de sacerdotibus autem Iedaia Io- 9:10 And of the priests; Jedaiah, and Jehoiarib, and Jachin,

iarib et Iachin 9:11 Azarias quoque filius Helciae filii Mo- 9:11 And Azariah the son of Hilkiah, the son of Meshul-

lam, the son of Zadok, the son of Meraioth, the son of Ahitub, the ruler of the house of God;

sollam filii Sadoc filii Maraioth filii Ahitob pontifex domus Dei 9:12 porro Adaias filius Hieroam filii Phasor filii Mel- 9:12 And Adaiah the son of Jeroham, the son of Pashur,

the son of Malchijah, and Maasiai the son of Adiel, the son of Jahzerah, the son of Meshullam, the son of

chia et Masaia filius Adihel filii Iezra filii Mosollam filii Meshillemith, the son of Immer;

Mosollamoth filii Emmer 9:13 fratres quoque eorum prin- 9:13 And their brethren, heads of the house of their fathers, a thousand and seven hundred and threescore; very able men for the work of the service of the house of God.

cipes per familias suas mille septingenti sexaginta fortissimi robore ad faciendum opus ministerii in domo Dei 9:14 de 9:14 And of the Levites; Shemaiah the son of Hasshub,

the son of Azrikam, the son of Hashabiah, of the sons of Merari;

Leuitis autem Semeia filius Assub filii Ezricam filii Asebiu de filiis Merari 9:15 Bacbacar quoque carpentarius et Ga- 9:15 And Bakbakkar, Heresh, and Galal, and Mattaniah

the son of Micah, the son of Zichri, the son of Asaph;

lal et Mathania filius Micha filii Zechri filii Asaph 9:16 et 9:16 And Obadiah the son of Shemaiah, the son of Galal,

the son of Jeduthun, and Berechiah the son of Asa, the son of Elkanah, that dwelt in the villages of the Netophathites.

Obdia filius Semeiae filii Galal filii Idithun et Barachia filius Asa filii Helcana qui habitauit in atriis Netophathi 9:17 9:17 And the porters were, Shallum, and Akkub, and Tal-

mon, and Ahiman, and their brethren: Shallum was the chief;

ianitores autem Sellum et Acub et Telmon et Ahiman et fra-

ter eorum Sellum princeps 9:18 usque ad illud tempus in 9:18 Who hitherto waited in the king’s gate eastward: they

were porters in the companies of the children of Levi.

porta Regis ad orientem obseruabant per uices suas de filiis Leui 9:19 Sellum uero filius Core filii Abiasaph filii Core 9:19 And Shallum the son of Kore, the son of Ebiasaph,

the son of Korah, and his brethren, of the house of his father, the Korahites, were over the work of the service, keepers of the gates of the tabernacle: and their fathers, being over the host of the LORD, were keepers of the entry.

cum fratribus suis et domo patris sui hii sunt Coritae super opera ministerii custodes uestibulorum tabernaculi et familiae eorum per uices castrorum Domini custodientes introitum 9:20 Finees autem filius Eleazar erat dux eorum co- 9:20 And Phinehas the son of Eleazar was the ruler over

them in time past, and the LORD was with him.

ram Domino 9:21 porro Zaccharias filius Mosollamia ia- 9:21 And Zechariah the son of Meshelemiah was porter of

the door of the tabernacle of the congregation.

nitor portae tabernaculi testimonii 9:22 omnes hii electi in 9:22 All these which were chosen to be porters in the gates

were two hundred and twelve. These were reckoned by their genealogy in their villages, whom David and Samuel

ostiarios per portas ducenti duodecim et descripti in uillis the seer did ordain in their set office. propriis quos constituerunt Dauid et Samuhel uidens in fide

sua 9:23 tam ipsos quam filios eorum in ostiis domus Do- 9:23 So they and their children had the oversight of the gates of the house of the LORD, namely, the house of the tabernacle, by wards.

Biblia Sacra Vulgata

9:24 In four quarters were the porters, toward the east, mini et in tabernaculo uicibus suis 9:24 per quattuor uentos

west, north, and south.

erant ostiarii id est ad orientem et ad occidentem ad aqui-

9:25 And their brethren, which were in their villages, were lonem et ad austrum 9:25 fratres autem eorum in uiculis to come after seven days from time to time with them.

morabantur et ueniebant in sabbatis suis de tempore usque

9:26 For these Levites, the four chief porters, were in their ad tempus 9:26 his quattuor Leuitis creditus erat omnis nuset office, and were over the chambers and treasuries of the house of God.

merus ianitorum et erant super exedras et thesauros domus

9:27 And they lodged round about the house of God, be- Domini 9:27 per gyrum quoque templi Domini morabantur

cause the charge was upon them, and the opening thereof every morning pertained to them.

in custodiis suis ut cum tempus fuisset ipsi mane aperirent

9:28 And certain of them had the charge of the ministering fores 9:28 de horum grege erant et super uasa ministerii ad

vessels, that they should bring them in and out by tale.

9:29 Some of them also were appointed to oversee the ves- numerum enim et inferebantur uasa et efferebantur 9:29 de

sels, and all the instruments of the sanctuary, and the fine flour, and the wine, and the oil, and the frankincense, and

the spices. ipsis et qui credita habebant utensilia sanctuarii praeerant 9:30 And some of the sons of the priests made the oint- similae et uino et oleo et turi et aromatibus 9:30 filii au-

ment of the spices. 9:31 And Mattithiah, one of the Levites, who was the first- tem sacerdotum unguenta ex aromatibus conficiebant 9:31 born of Shallum the Korahite, had the set office over the things that were made in the pans.

et Matthathias Leuites primogenitus Sellum Coritae praefec-

9:32 And other of their brethren, of the sons of the Ko- tus erat eorum quae in sartagine frigebantur 9:32 porro de

hathites, were over the shewbread, to prepare it every sabbath.

filiis Caath fratribus eorum super panes erant propositionis

9:33 And these are the singers, chief of the fathers of the ut semper nouos per singula sabbata praepararent 9:33 hii

Levites, who remaining in the chambers were free: for they were employed in that work day and night.

sunt principes cantorum per familias Leuitarum qui in exedris morabantur ita ut die et nocte iugiter suo ministerio de-

9:34 These chief fathers of the Levites were chief through- seruirent 9:34 capita Leuitarum per familias suas principes

out their generations; these dwelt at Jerusalem.

9:35 And in Gibeon dwelt the father of Gibeon, Jehiel, manserunt in Hierusalem 9:35 in Gabaon autem commorati

whose wife’s name was Maachah:

9:36 And his firstborn son Abdon, then Zur, and Kish, and sunt pater Gabaon Iaihel et nomen uxoris eius Maacha 9:36

Baal, and Ner, and Nadab.

filius primogenitus eius Abdon et Sur et Cis et Baal et Ner et

9:37 And Gedor, and Ahio, and Zechariah, and Mikloth. Nadab 9:37 Gedor quoque et Ahio et Zaccharias et Macel-9:38 And Mikloth begat Shimeam. And they also loth 9:38 porro Macelloth genuit Semmaam isti habitaue-

dwelt with their brethren at Jerusalem, over against their brethren.

runt e regione fratrum suorum in Hierusalem cum fratribus

9:39 And Ner begat Kish; and Kish begat Saul; and Saul suis 9:39 Ner autem genuit Cis et Cis genuit Saul et Saul

begat Jonathan, and Malchishua, and Abinadab, and Eshbaal.

genuit Ionathan et Melchisuae et Abinadab et Esbaal 9:40

9:40 And the son of Jonathan was Meribbaal: and Meribbaal begat Micah.

filius autem Ionathan Meribbaal et Meribbaal genuit Micha

9:41 And the sons of Micah were, Pithon, and Melech, and 9:41 porro filii Micha Phiton et Malech et Thara 9:42 Ahaz

Tahrea, and Ahaz. 9:42 And Ahaz begat Jarah; and Jarah begat Alemeth, and Azmaveth, and Zimri; and Zimri begat Moza;

autem genuit Iara et Iara genuit Alamath et Azmoth et Zamri

9:43 And Moza begat Binea; and Rephaiah his son, et Zamri genuit Mosa 9:43 Mosa uero genuit Baana cuius

Eleasah his son, Azel his son.

9:44 And Azel had six sons, whose names are these, filius Raphaia genuit Elasa de quo ortus est Esel 9:44 porro

Azrikam, Bocheru, and Ishmael, and Sheariah, and Obadiah, and Hanan: these were the sons of Azel.

Liber I Paralipomenon

Esel sex filios habuit his nominibus Ezricam Bochru Ismahel Saria Obdia Anan hii filii Esel

10:1 Philisthim autem pugnabant contra Israhel fugerunt- 10:1 Now the Philistines fought against Israel; and the

men of Israel fled from before the Philistines, and fell down slain in mount Gilboa.

que uiri Israhel Palestinos et ceciderunt uulnerati in monte

Gelboe 10:2 cumque adpropinquassent Philisthei perse- 10:2 And the Philistines followed hard after Saul, and after his sons; and the Philistines slew Jonathan, and Abinadab, and Malchishua, the sons of Saul.

quentes Saul et filios eius percusserunt Ionathan et Abina-

dab et Melchisuae filios Saul 10:3 et adgrauatum est proe- 10:3 And the battle went sore against Saul, and the archers

hit him, and he was wounded of the archers.

lium contra Saul inueneruntque eum sagittarii et uulnerauerunt iaculis 10:4 et dixit Saul ad armigerum suum euagina 10:4 Then said Saul to his armourbearer, Draw thy sword,

and thrust me through therewith; lest these uncircumcised come and abuse me. But his armourbearer would not; for

gladium tuum et interfice me ne forte ueniant incircumcisi he was sore afraid. So Saul took a sword, and fell upon it. isti et inludant mihi noluit autem armiger eius hoc facere timore perterritus arripuit igitur Saul ensem et inruit in eum 10:5 quod cum uidisset armiger eius uidelicet mortuum esse 10:5 And when his armourbearer saw that Saul was dead,

he fell likewise on the sword, and died.

Saul inruit etiam ipse in gladium suum et mortuus est 10:6 10:6 So Saul died, and his three sons, and all his house

died together.

interiit ergo Saul et tres filii eius et omnis domus illius pariter concidit 10:7 quod cum uidissent uiri Israhel qui ha- 10:7 And when all the men of Israel that were in the valley

saw that they fled, and that Saul and his sons were dead, then they forsook their cities, and fled: and the Philistines

bitabant in campestribus fugerunt et Saul ac filiis eius mor- came and dwelt in them. tuis dereliquerunt urbes suas et huc illucque dispersi sunt ueneruntque Philisthim et habitauerunt in eis 10:8 die igi- 10:8 And it came to pass on the morrow, when the

Philistines came to strip the slain, that they found Saul and his sons fallen in mount Gilboa.

tur altero detrahentes Philisthim spolia caesorum inuenerunt Saul et filios eius iacentes in monte Gelboe 10:9 cumque 10:9 And when they had stripped him, they took his head,

and his armour, and sent into the land of the Philistines round about, to carry tidings unto their idols, and to the

spoliassent eum et amputassent caput armisque nudassent people. miserunt in terram suam ut circumferretur et ostenderetur idolorum templis et populis 10:10 arma autem eius consec- 10:10 And they put his armour in the house of their gods,

and fastened his head in the temple of Dagon.

rauerunt in fano dei sui et caput adfixerunt in templo Dagon 10:11 hoc cum audissent uiri Iabesgalaad omnia scilicet quae 10:11 And when all Jabeshgilead heard all that the

Philistines had done to Saul,

Philisthim fecerunt super Saul 10:12 consurrexerunt singuli 10:12 They arose, all the valiant men, and took away the

body of Saul, and the bodies of his sons, and brought them to Jabesh, and buried their bones under the oak in Jabesh,

uirorum fortium et tulerunt cadauera Saul et filiorum eius and fasted seven days. adtuleruntque ea in Iabes et sepelierunt ossa eorum subter quercum quae erat in Iabes et ieiunauerunt septem diebus 10:13 mortuus est ergo Saul propter iniquitates suas eo quod 10:13 So Saul died for his transgression which he commit-

ted against the LORD, even against the word of the LORD, which he kept not, and also for asking counsel of one that

praeuaricatus sit mandatum Domini quod praeceperat et non had a familiar spirit, to enquire of it; custodierit illud sed insuper etiam pythonissam consuluerit

Biblia Sacra Vulgata

10:14 And enquired not of the LORD: therefore he slew 10:14 nec sperauerit in Domino propter quod et interfecit eum

him, and turned the kingdom unto David the son of Jesse.

et transtulit regnum eius ad Dauid filium Isai

11:1 Then all Israel gathered themselves to David unto He- 11:1 congregatus est igitur omnis Israhel ad Dauid in Hebron, saying, Behold, we are thy bone and thy flesh. 11:2 And moreover in time past, even when Saul was king, bron dicens os tuum sumus et caro tua 11:2 heri quoque et thou wast he that leddest out and broughtest in Israel: and the LORD thy God said unto thee, Thou shalt feed my people Israel, and thou shalt be ruler over my people Israel.

nudius tertius cum adhuc regnaret Saul tu eras qui educebas et introducebas Israhel tibi enim dixit Dominus Deus tuus tu pasces populum meum Israhel et tu eris princeps super eum

11:3 Therefore came all the elders of Israel to the king to 11:3 uenerunt ergo omnes maiores natu Israhel ad regem in Hebron; and David made a covenant with them in Hebron before the LORD; and they anointed David king over Is-

rael, according to the word of the LORD by Samuel. Hebron et iniit Dauid cum eis foedus coram Domino unxe-

runtque eum regem super Israhel iuxta sermonem Domini

11:4 And David and all Israel went to Jerusalem, which quem locutus est in manu Samuhel 11:4 abiit quoque Dais Jebus; where the Jebusites were, the inhabitants of the land.

uid et omnis Israhel in Hierusalem haec est Iebus ubi erant

11:5 And the inhabitants of Jebus said to David, Thou shalt Iebusei habitatores terrae 11:5 dixeruntque qui habitabant

not come hither. Nevertheless David took the castle of Zion, which is the city of David.

in Iebus ad Dauid non ingredieris huc porro Dauid cepit ar-

11:6 And David said, Whosoever smiteth the Jebusites cem Sion quae est ciuitas Dauid 11:6 dixitque omnis qui

first shall be chief and captain. So Joab the son of Zeruiah went first up, and was chief.

percusserit Iebuseum in primis erit princeps et dux ascendit

11:7 And David dwelt in the castle; therefore they called igitur primus Ioab filius Saruiae et factus est princeps 11:7

it the city of David.

habitauit autem Dauid in arce et idcirco appellata est ciuitas

11:8 And he built the city round about, even from Millo Dauid 11:8 aedificauitque urbem in circuitu a Mello usque

round about: and Joab repaired the rest of the city.

11:9 So David waxed greater and greater: for the LORD ad gyrum Ioab autem reliqua urbis extruxit 11:9 proficie-

of hosts was with him.

batque Dauid uadens et crescens et Dominus exercituum erat

11:10 These also are the chief of the mighty men whom cum eo 11:10 hii principes uirorum fortium Dauid qui adiu-

David had, who strengthened themselves with him in his kingdom, and with all Israel, to make him king, according

to the word of the LORD concerning Israel. uerunt eum ut rex fieret super omnem Israhel iuxta uerbum 11:11 And this is the number of the mighty men whom Domini quod locutus est ad Israhel 11:11 et iste numerus

David had; Jashobeam, an Hachmonite, the chief of the captains: he lifted up his spear against three hundred slain

by him at one time. robustorum Dauid Iesbaam filius Achamoni princeps inter

triginta iste leuauit hastam suam super trecentos uulneratos

11:12 And after him was Eleazar the son of Dodo, the una uice 11:12 et post eum Eleazar filius patrui eius Ahoi-

Ahohite, who was one of the three mighties.

11:13 He was with David at Pasdammim, and there the tes qui erat inter tres potentes 11:13 iste fuit cum Dauid in

Philistines were gathered together to battle, where was a parcel of ground full of barley; and the people fled from

before the Philistines. Aphesdommim quando Philisthim congregati sunt ad locum

illum in proelium et erat ager regionis illius plenus hordeo

11:14 And they set themselves in the midst of that parcel, fugeratque populus a facie Philisthinorum 11:14 hic stetit in

and delivered it, and slew the Philistines; and the LORD saved them by a great deliverance.

medio agri et defendit eum cumque percussisset Philistheos

11:15 Now three of the thirty captains went down to the dedit Dominus salutem magnam populo suo 11:15 descen-

rock to David, into the cave of Adullam; and the host of the Philistines encamped in the valley of Rephaim.

Liber I Paralipomenon

derunt autem tres de triginta principibus ad petram in qua erat Dauid ad speluncam Odollam quando Philisthim fuerant

castrametati in ualle Raphaim 11:16 porro Dauid erat in 11:16 And David was then in the hold, and the Philistines’

garrison was then at Bethlehem.

praesidio et statio Philisthinorum in Bethleem 11:17 desi- 11:17 And David longed, and said, Oh that one would give

me drink of the water of the well of Bethlehem, that is at the gate!

derauit igitur Dauid et dixit o si quis daret mihi aquam de cisterna Bethleem quae est in porta 11:18 tres ergo isti per me- 11:18 And the three brake through the host of the

Philistines, and drew water out of the well of Bethlehem, that was by the gate, and took it, and brought it to David: but David would not drink of it, but poured it out to the LORD.

dia castra Philisthinorum perrexerunt et hauserunt aquam de cisterna Bethleem quae erat in porta et adtulerunt ad Dauid ut biberet qui noluit sed magis libauit illam Domino 11:19 11:19 And said, My God forbid it me, that I should do this

thing: shall I drink the blood of these men that have put their lives in jeopardy? for with the jeopardy of their lives they brought it. Therefore he would not drink it. These things did these three mightiest.

dicens absit ut in conspectu Dei mei hoc faciam et sanguinem uirorum istorum bibam quia in periculo animarum suarum adtulerunt mihi aquam et ob hanc causam noluit bibere haec fecerunt tres robustissimi 11:20 Abisai quoque frater 11:20 And Abishai the brother of Joab, he was chief of

the three: for lifting up his spear against three hundred, he slew them, and had a name among the three.

Ioab ipse erat princeps trium et ipse leuauit hastam suam contra trecentos uulneratos et ipse erat inter tres nominatissimus 11:21 inter tres secundos inclitus et princeps eorum 11:21 Of the three, he was more honourable than the two;

for he was their captain: howbeit he attained not to the first three.

uerumtamen usque ad tres primos non peruenerat 11:22 Ba-

11:22 Benaiah the son of Jehoiada, the son of a valiant man of Kabzeel, who had done many acts; he slew two lionlike men of Moab: also he went down and slew a lion in a pit in a snowy day.

naia filius Ioiadae uiri robustissimi qui multa opera perpetrarat de Capsehel ipse percussit duos Arihel Moab et ipse descendit et interfecit leonem in media cisterna tempore niuis 11:23 et ipse percussit uirum aegyptium cuius statura erat 11:23 And he slew an Egyptian, a man of great stature, five

cubits high; and in the Egyptian’s hand was a spear like a weaver’s beam; and he went down to him with a staff, and plucked the spear out of the Egyptian’s hand, and slew him with his own spear.

quinque cubitorum et habebat lanceam ut liciatorium texentium descendit ergo ad eum cum uirga et rapuit hastam quam tenebat manu et interfecit eum hasta sua 11:24 haec fecit 11:24 These things did Benaiah the son of Jehoiada, and

had the name among the three mighties.

Banaia filius Ioiada qui erat inter tres robustos nominatissimus 11:25 inter triginta primus uerumtamen ad tres usque 11:25 Behold, he was honourable among the thirty, but

attained not to the first three: and David set him over his guard.

non peruenerat posuit autem eum Dauid ad auriculam suam 11:26 porro fortissimi in exercitu Asahel frater Ioab et Elea- 11:26 Also the valiant men of the armies were, Asahel the

brother of Joab, Elhanan the son of Dodo of Bethlehem,

nan filius patrui eius de Bethleem 11:27 Semmoth Arorites 11:27 Shammoth the Harorite, Helez the Pelonite, Helles Phallonites 11:28 Iras filius Acces Thecuites Abie- 11:28 Ira the son of Ikkesh the Tekoite, Abiezer the An-

tothite,

zer Anathothites 11:29 Sobbochai Asothites Ilai Ahoites 11:29 Sibbecai the Hushathite, Ilai the Ahohite, 11:30 Marai Netophathites Heled filius Baana Netophathites 11:30 Maharai the Netophathite, Heled the son of Baanah

the Netophathite,

11:31 Ethai filius Ribai de Gabaath filiorum Beniamin Banaia 11:31 Ithai the son of Ribai of Gibeah, that pertained to

the children of Benjamin, Benaiah the Pirathonite,

Biblia Sacra Vulgata

11:32 Hurai of the brooks of Gaash, Abiel the Arbathite, Pharathonites 11:32 Vri de torrente Gaas Abial Arabathi-

11:33 Azmaveth the Baharumite, Eliahba the Shaalbonite, tes Azmoth Bauramites Eliaba Salabonites 11:33 filii Asom

11:34 The sons of Hashem the Gizonite, Jonathan the son Gezonites Ionathan filius Sega Ararites 11:34 Ahiam filius

of Shage the Hararite, 11:35 Ahiam the son of Sacar the Hararite, Eliphal the son Sachar Ararites 11:35 Eliphal filius Vr 11:36 Apher Meof Ur, 11:36 Hepher the Mecherathite, Ahijah the Pelonite,

churathites Ahia Phellonites 11:37 Asrai Carmelites Noo-

11:37 Hezro the Carmelite, Naarai the son of Ezbai,

11:38 Joel the brother of Nathan, Mibhar the son of Hag- rai filius Azbi 11:38 Iohel frater Nathan Mabar filius Agarai

geri,

11:39 Zelek the Ammonite, Naharai the Berothite, the ar- 11:39 Sellec Ammonites Noorai Berothites armiger Ioab filii

mourbearer of Joab the son of Zeruiah,

11:40 Ira the Ithrite, Gareb the Ithrite, Saruiae 11:40 Iras Iethreus Gareb Iethreus 11:41 Vrias Et-

11:41 Uriah the Hittite, Zabad the son of Ahlai,

11:42 Adina the son of Shiza the Reubenite, a captain of theus Zabad filius Ooli 11:42 Adina filius Seza Rubenites

the Reubenites, and thirty with him,

11:43 Hanan the son of Maachah, and Joshaphat the Mith- princeps Rubenitarum et cum eo triginta 11:43 Hanan filius

nite,

11:44 Uzzia the Ashterathite, Shama and Jehiel the sons Maacha et Iosaphat Mathanites 11:44 Ozias Astharothites

of Hothan the Aroerite, 11:45 Jediael the son of Shimri, and Joha his brother, the Semma et Iaihel filii Hotam Aroerites 11:45 Iedihel filius Tizite, 11:46 Eliel the Mahavite, and Jeribai, and Joshaviah, the Samri et Ioha frater eius Thosaites 11:46 Elihel Maumites sons of Elnaam, and Ithmah the Moabite, Eliel, and Obed, and Jasiel the Mesobaite.

et Ieribai et Iosaia filii Elnaem et Iethma Moabites Elihel et Obed et Iasihel de Masobia

12:1 Now these are they that came to David to Ziklag, 12:1 hii quoque uenerunt ad Dauid in Siceleg cum adhuc

while he yet kept himself close because of Saul the son of Kish: and they were among the mighty men, helpers of the

war. fugeret Saul filium Cis qui erant fortissimi et egregii pug-12:2 They were armed with bows, and could use both the natores 12:2 tendentes arcum et utraque manu fundis saxa

right hand and the left in hurling stones and shooting arrows out of a bow, even of Saul’s brethren of Benjamin.

iacientes et dirigentes sagittas de fratribus Saul ex Beniamin

12:3 The chief was Ahiezer, then Joash, the sons of She- 12:3 princeps Ahiezer et Ioas filii Sammaa Gabathites et Ia-

maah the Gibeathite; and Jeziel, and Pelet, the sons of Azmaveth; and Berachah, and Jehu the Antothite.

zihel et Phallet filii Azmoth et Baracha et Ieu Anathothites

12:4 And Ismaiah the Gibeonite, a mighty man among the 12:4 Samaias quoque Gabaonites fortissimus inter triginta et

thirty, and over the thirty; and Jeremiah, and Jahaziel, and Johanan, and Josabad the Gederathite,

super triginta Hieremias et Iezihel et Iohanan et Iezbad Ga-

12:5 Eluzai, and Jerimoth, and Bealiah, and Shemariah, derothites 12:5 Eluzai et Ierimuth et Baalia et Samaria et

and Shephatiah the Haruphite,

12:6 Elkanah, and Jesiah, and Azareel, and Joezer, and Saphatia Aruphites 12:6 Helcana et Iesia et Azrahel et Ioe-

Jashobeam, the Korhites,

12:7 And Joelah, and Zebadiah, the sons of Jeroham of zer et Iesbaam de Careim 12:7 Ioeela quoque et Zabadia

Gedor.

12:8 And of the Gadites there separated themselves unto filii Ieroam de Gedor 12:8 sed et de Gaddi transfugerunt ad

David into the hold to the wilderness men of might, and men of war fit for the battle, that could handle shield and buckler, whose faces were like the faces of lions, and were as swift as the roes upon the mountains;

Dauid cum lateret in deserto uiri robustissimi et pugnatores optimi tenentes clypeum et hastam facies eorum quasi facies

12:9 Ezer the first, Obadiah the second, Eliab the third, leonis et ueloces quasi capreae in montibus 12:9 Ezer prin-12:10 Mishmannah the fourth, Jeremiah the fifth, ceps Obdias secundus Eliab tertius 12:10 Masmana quar-12:11 Attai the sixth, Eliel the seventh, tus Hieremias quintus 12:11 Hetthi sextus Helihel septimus 12:12 Johanan the eighth, Elzabad the ninth, 12:12 Iohanan octauus Helzebad nonus 12:13 Hieremias de-

12:13 Jeremiah the tenth, Machbanai the eleventh.

Liber I Paralipomenon

cimus Bachannai undecimus 12:14 hii de filiis Gad princi- 12:14 These were of the sons of Gad, captains of the host:

one of the least was over an hundred, and the greatest over a thousand.

pes exercitus nouissimus centum militibus praeerat et maximus mille 12:15 isti sunt qui transierunt Iordanem mense 12:15 These are they that went over Jordan in the first

month, when it had overflown all his banks; and they put to flight all them of the valleys, both toward the east, and

primo quando inundare consueuit super ripas suas et omnes toward the west. fugauerunt qui morabantur in uallibus ad orientalem plagam et occidentalem 12:16 uenerunt autem et de Beniamin et 12:16 And there came of the children of Benjamin and

Judah to the hold unto David.

de Iuda ad praesidium in quo morabatur Dauid 12:17 eg- 12:17 And David went out to meet them, and answered

and said unto them, If ye be come peaceably unto me to help me, mine heart shall be knit unto you: but if ye be come to betray me to mine enemies, seeing there is no wrong in mine hands, the God of our fathers look thereon, and rebuke it.

ressusque est Dauid obuiam eis et ait si pacifice uenistis ad me ut auxiliemini mihi cor meum iungatur uobis si autem insidiamini mihi pro aduersariis meis cum ego iniquitatem in manibus non habeam uideat Deus patrum nostrorum et iudicet 12:18 spiritus uero induit Amessai principem inter 12:18 Then the spirit came upon Amasai, who was chief

of the captains, and he said, Thine are we, David, and on thy side, thou son of Jesse: peace, peace be unto thee, and peace be to thine helpers; for thy God helpeth thee. Then David received them, and made them captains of the band.

triginta et ait tui sumus o Dauid et tecum fili Isai pax pax tibi et pax adiutoribus tuis te enim adiuuat Deus tuus suscepit ergo eos Dauid et constituit principes turmae 12:19 12:19 And there fell some of Manasseh to David, when he

came with the Philistines against Saul to battle: but they helped them not: for the lords of the Philistines upon advisement sent him away, saying, He will fall to his master Saul to the jeopardy of our heads.

porro de Manasse transfugerunt ad Dauid quando ueniebat cum Philisthim aduersum Saul ut pugnaret et non dimicauit cum eis quia inito consilio remiserunt eum principes Philisthinorum dicentes periculo capitis nostri reuertetur ad dominum suum Saul 12:20 quando igitur reuersus est in Siceleg 12:20 As he went to Ziklag, there fell to him of Manasseh,

Adnah, and Jozabad, and Jediael, and Michael, and Jozabad, and Elihu, and Zilthai, captains of the thousands that

transfugerunt ad eum de Manasse Ednas et Iozabad et Iedi- were of Manasseh. hel et Michahel et Iozabad et Heliu et Salathi principes milium in Manasse 12:21 hii praebuerunt auxilium Dauid ad- 12:21 And they helped David against the band of the

rovers: for they were all mighty men of valour, and were captains in the host.

uersum latrunculos omnes enim erant uiri fortissimi et facti sunt principes in exercitu 12:22 sed et per singulos dies ue- 12:22 For at that time day by day there came to David to

help him, until it was a great host, like the host of God.

niebant ad Dauid ad auxiliandum ei usque dum fieret grandis numerus quasi exercitus Dei 12:23 iste quoque est numerus 12:23 And these are the numbers of the bands that were

ready armed to the war, and came to David to Hebron, to turn the kingdom of Saul to him, according to the word of

principum exercitus qui uenerunt ad Dauid cum esset in He- the LORD. bron ut transferrent regnum Saul ad eum iuxta uerbum Domini 12:24 filii Iuda portantes clypeum et hastam sex mi- 12:24 The children of Judah that bare shield and spear

were six thousand and eight hundred, ready armed to the war.

lia octingenti expediti ad proelium 12:25 de filiis Symeon

12:25 Of the children of Simeon, mighty men of valour for the war, seven thousand and one hundred.

uirorum fortissimorum ad pugnandum septem milia centum 12:26 de filiis Leui quattuor milia sescenti 12:27 Ioiada quo- 12:26 Of the children of Levi four thousand and six hun-

dred.

12:27 And Jehoiada was the leader of the Aaronites, and with him were three thousand and seven hundred;

Biblia Sacra Vulgata

que princeps de stirpe Aaron et cum eo tria milia septingenti

12:28 And Zadok, a young man mighty of valour, and of 12:28 Sadoc etiam puer egregiae indolis et domus patris eius

his father’s house twenty and two captains.

12:29 And of the children of Benjamin, the kindred of principes uiginti duo 12:29 de filiis autem Beniamin fratri-

Saul, three thousand: for hitherto the greatest part of them had kept the ward of the house of Saul.

bus Saul tria milia magna enim pars eorum adhuc sequeba-

12:30 And of the children of Ephraim twenty thousand and tur domum Saul 12:30 porro de filiis Ephraim uiginti milia

eight hundred, mighty men of valour, famous throughout the house of their fathers.

octingenti fortissimi robore uiri nominati in cognationibus

12:31 And of the half tribe of Manasseh eighteen thou- suis 12:31 et ex dimidia parte tribus Manasse decem et octo

sand, which were expressed by name, to come and make David king.

milia singuli per nomina sua uenerunt ut constituerent re-

12:32 And of the children of Issachar, which were men gem Dauid 12:32 de filiis quoque Isachar uiri eruditi qui no-

that had understanding of the times, to know what Israel ought to do; the heads of them were two hundred; and all

their brethren were at their commandment. rant singula tempora ad praecipiendum quid facere deberet

Israhel principes ducenti omnis autem reliqua tribus eorum

12:33 Of Zebulun, such as went forth to battle, expert in consilium sequebatur 12:33 porro de Zabulon qui egredie-

war, with all instruments of war, fifty thousand, which could keep rank: they were not of double heart.

bantur ad proelium et stabant in acie instructi armis bellicis quinquaginta milia uenerunt in auxilium non in corde duplici

12:34 And of Naphtali a thousand captains, and with them 12:34 et de Nepthali principes mille et cum eis instructa cly-

with shield and spear thirty and seven thousand.

12:35 And of the Danites expert in war twenty and eight peo et hasta triginta septem milia 12:35 de Dan etiam prae-

thousand and six hundred.

12:36 And of Asher, such as went forth to battle, expert in parata ad proelium uiginti octo milia sescentorum 12:36 et

war, forty thousand.

de Aser egredientes ad pugnam et in acie prouocantes qua-

12:37 And on the other side of Jordan, of the Reubenites, draginta milia 12:37 trans Iordanem autem de filiis Ruben

and the Gadites, and of the half tribe of Manasseh, with all manner of instruments of war for the battle, an hundred

and twenty thousand. et Gad et dimidia parte tribus Manasse instructa armis bel-12:38 All these men of war, that could keep rank, came licis centum uiginti milia 12:38 omnes isti uiri bellatores et

with a perfect heart to Hebron, to make David king over all Israel: and all the rest also of Israel were of one heart

to make David king. expediti ad pugnandum corde perfecto uenerunt in Hebron ut constituerent regem Dauid super uniuersum Israhel sed et omnes reliqui ex Israhel uno corde erant ut rex fieret Dauid

12:39 And there they were with David three days, eating 12:39 fueruntque ibi apud Dauid tribus diebus comedentes

and drinking: for their brethren had prepared for them.

12:40 Moreover they that were nigh them, even unto Is- et bibentes praeparauerunt enim eis fratres sui 12:40 sed et

sachar and Zebulun and Naphtali, brought bread on asses, and on camels, and on mules, and on oxen, and meat, meal, cakes of figs, and bunches of raisins, and wine, and oil, and oxen, and sheep abundantly: for there was joy in Israel.

qui iuxta eos erant usque ad Isachar et Zabulon et Nepthalim adferebant panes in asinis et camelis et mulis et bubus ad uescendum farinam palatas uuam passam uinum oleum boues arietes ad omnem copiam gaudium quippe erat in Israhel

13:1 And David consulted with the captains of thousands 13:1 iniit autem consilium Dauid cum tribunis et centuand hundreds, and with every leader.

13:2 And David said unto all the congregation of Israel, rionibus et uniuersis principibus 13:2 et ait ad omnem coe-

If it seem good unto you, and that it be of the LORD our God, let us send abroad unto our brethren every where, that are left in all the land of Israel, and with them also to the priests and Levites which are in their cities and suburbs, that they may gather themselves unto us:

tum Israhel si placet uobis et a Domino Deo nostro egredi-

Liber I Paralipomenon

tur sermo quem loquor mittamus ad fratres nostros reliquos in uniuersas regiones Israhel et ad sacerdotes et Leuitas qui habitant in suburbanis urbium ut congregentur ad nos 13:3 13:3 And let us bring again the ark of our God to us: for

we enquired not at it in the days of Saul.

et reducamus arcam Dei nostri ad nos non enim requisiuimus eam in diebus Saul 13:4 et respondit uniuersa multi- 13:4 And all the congregation said that they would do so:

for the thing was right in the eyes of all the people.

tudo ut ita fieret placuerat enim sermo omni populo 13:5 13:5 So David gathered all Israel together, from Shihor of Egypt even unto the entering of Hemath, to bring the ark of God from Kirjathjearim.

congregauit ergo Dauid cunctum Israhel a Sior Aegypti usque dum ingrediaris Emath ut adduceret arcam Dei de Cariathiarim 13:6 et ascendit Dauid et omnis uir Israhel ad 13:6 And David went up, and all Israel, to Baalah, that

is, to Kirjathjearim, which belonged to Judah, to bring up thence the ark of God the LORD, that dwelleth between

collem Cariathiarim quae est in Iuda ut adferrent inde arcam the cherubims, whose name is called on it. Dei Domini sedentis super cherubin ubi inuocatum est nomen eius 13:7 inposueruntque arcam Dei super plaustrum 13:7 And they carried the ark of God in a new cart out of

the house of Abinadab: and Uzza and Ahio drave the cart.

nouum de domo Aminadab Oza autem et fratres eius minabant plaustrum 13:8 porro Dauid et uniuersus Israhel lu- 13:8 And David and all Israel played before God with all

their might, and with singing, and with harps, and with psalteries, and with timbrels, and with cymbals, and with

debant coram Deo omni uirtute in canticis et in citharis et trumpets. psalteriis et tympanis et cymbalis et tubis 13:9 cum autem 13:9 And when they came unto the threshingfloor of Chi-

don, Uzza put forth his hand to hold the ark; for the oxen stumbled.

peruenissent ad aream Chidon tetendit Oza manum suam ut sustentaret arcam bos quippe lasciuiens paululum inclinauerat eam 13:10 iratus est itaque Dominus contra Ozam et 13:10 And the anger of the LORD was kindled against

Uzza, and he smote him, because he put his hand to the ark: and there he died before God.

percussit eum eo quod contigisset arcam et mortuus est ibi coram Deo 13:11 contristatusque Dauid eo quod diuisisset 13:11 And David was displeased, because the LORD had

made a breach upon Uzza: wherefore that place is called Perezuzza to this day.

Dominus Ozam uocauit locum illum Diuisio Oza usque in praesentem diem 13:12 et timuit Deum tunc temporis di- 13:12 And David was afraid of God that day, saying, How

shall I bring the ark of God home to me?

cens quomodo possum ad me introducere arcam Dei 13:13 13:13 So David brought not the ark home to himself to

the city of David, but carried it aside into the house of Obededom the Gittite.

et ob hanc causam non eam adduxit ad se hoc est in ciuitatem Dauid sed auertit in domum Obededom Getthei 13:14 13:14 And the ark of God remained with the family of

Obededom in his house three months. And the LORD blessed the house of Obededom, and all that he had.

mansit ergo arca Dei in domo Obededom tribus mensibus et benedixit Dominus domui eius et omnibus quae habebat

14:1 misit quoque Hiram rex Tyri nuntios ad Dauid et ligna 14:1 Now Hiram king of Tyre sent messengers to David,

and timber of cedars, with masons and carpenters, to build him an house.

cedrina et artifices parietum lignorumque ut aedificarent ei domum 14:2 cognouitque Dauid eo quod confirmasset eum 14:2 And David perceived that the LORD had confirmed

him king over Israel, for his kingdom was lifted up on high, because of his people Israel.

Dominus in regem super Israhel et subleuatum esset regnum suum super populum eius Israhel 14:3 accepit quoque Da- 14:3 And David took more wives at Jerusalem: and David

begat more sons and daughters.

Biblia Sacra Vulgata

14:4 Now these are the names of his children which he uid alias uxores in Hierusalem genuitque filios et filias 14:4

had in Jerusalem; Shammua, and Shobab, Nathan, and Solomon,

et haec nomina eorum qui nati sunt ei in Hierusalem Sammu

14:5 And Ibhar, and Elishua, and Elpalet, et Sobab Nathan et Salomon 14:5 Iebar et Helisu et Eli-14:6 And Nogah, and Nepheg, and Japhia, phaleth 14:6 Noga quoque et Napheg et Iaphiae 14:7 Eli-

14:7 And Elishama, and Beeliada, and Eliphalet.

14:8 And when the Philistines heard that David was sama et Baliada et Eliphaleth 14:8 audientes autem Philis-

anointed king over all Israel, all the Philistines went up to seek David. And David heard of it, and went out against

them. thim eo quod unctus esset Dauid in regem super uniuersum

Israhel ascenderunt omnes ut quaererent eum quod cum au-

14:9 And the Philistines came and spread themselves in disset Dauid egressus est obuiam eis 14:9 porro Philisthim

the valley of Rephaim.

14:10 And David enquired of God, saying, Shall I go up uenientes diffusi sunt in ualle Raphaim 14:10 consuluitque

against the Philistines? And wilt thou deliver them into mine hand? And the LORD said unto him, Go up; for I

will deliver them into thine hand. Dauid Deum dicens si ascendam ad Philistheos si trades eos

in manu mea et dixit ei Dominus ascende et tradam eos in

14:11 So they came up to Baalperazim; and David smote manu tua 14:11 cumque illi ascendissent in Baalpharasim

them there. Then David said, God hath broken in upon mine enemies by mine hand like the breaking forth of waters: therefore they called the name of that place Baalperazim.

percussit eos ibi Dauid et dixit diuisit Deus inimicos meos per manum meam sicuti diuiduntur aquae et idcirco uocatum

14:12 And when they had left their gods there, David gave est nomen loci illius Baalpharasim 14:12 dereliqueruntque

a commandment, and they were burned with fire.

14:13 And the Philistines yet again spread themselves ibi deos suos quos Dauid iussit exuri 14:13 alia etiam uice

abroad in the valley.

14:14 Therefore David enquired again of God; and God Philisthim inruerunt et diffusi sunt in ualle 14:14 consuluit-

said unto him, Go not up after them; turn away from them, and come upon them over against the mulberry trees.

que rursum Dauid Deum et dixit ei Deus non ascendas post eos recede ab eis et uenies contra illos ex aduerso pirorum

14:15 And it shall be, when thou shalt hear a sound of 14:15 cumque audieris sonitum gradientis in cacumine piro-

going in the tops of the mulberry trees, that then thou shalt go out to battle: for God is gone forth before thee to smite

the host of the Philistines. rum tunc egredieris ad bellum egressus est enim Deus ante 14:16 David therefore did as God commanded him: and te ut percutiat castra Philisthim 14:16 fecit ergo Dauid sic-

they smote the host of the Philistines from Gibeon even to Gazer.

ut praeceperat ei Deus et percussit castra Philisthinorum de

14:17 And the fame of David went out into all lands; and Gabaon usque Gazera 14:17 diuulgatumque est nomen Da-

the LORD brought the fear of him upon all nations.

uid in uniuersis regionibus et Dominus dedit pauorem eius super omnes gentes

15:1 And David made him houses in the city of David, 15:1 fecit quoque sibi domos in ciuitate Dauid et aedifica-

and prepared a place for the ark of God, and pitched for it a tent.

uit locum arcae Dei tetenditque ei tabernaculum 15:2 tunc

15:2 Then David said, None ought to carry the ark of God but the Levites: for them hath the LORD chosen to carry the ark of God, and to minister unto him for ever.

dixit Dauid inlicitum est ut a quocumque portetur arca Dei nisi a Leuitis quos elegit Dominus ad portandum eam et ad

15:3 And David gathered all Israel together to Jerusalem, ministrandum sibi usque in aeternum 15:3 congregauitque to bring up the ark of the LORD unto his place, which he had prepared for it.

uniuersum Israhel in Hierusalem ut adferretur arca Dei in lo-

15:4 And David assembled the children of Aaron, and the cum suum quem praeparauerat ei 15:4 necnon et filios Aa-Levites:

Liber I Paralipomenon

ron et Leuitas 15:5 de filiis Caath Vrihel princeps fuit et fra- 15:5 Of the sons of Kohath; Uriel the chief, and his

brethren an hundred and twenty:

tres eius centum uiginti 15:6 de filiis Merari Asaia princeps 15:6 Of the sons of Merari; Asaiah the chief, and his

brethren two hundred and twenty:

et fratres eius ducenti uiginti 15:7 de filiis Gersom Iohel 15:7 Of the sons of Gershom; Joel the chief and his

brethren an hundred and thirty:

princeps et fratres eius centum triginta 15:8 de filiis Elisa- 15:8 Of the sons of Elizaphan; Shemaiah the chief, and his

brethren two hundred:

phan Semeias princeps et fratres eius ducenti 15:9 de filiis 15:9 Of the sons of Hebron; Eliel the chief, and his

brethren fourscore:

Hebron Elihel princeps et fratres eius octoginta 15:10 de fi- 15:10 Of the sons of Uzziel; Amminadab the chief, and

his brethren an hundred and twelve.

liis Ozihel Aminadab princeps et fratres eius centum duodecim 15:11 uocauitque Dauid Sadoc et Abiathar sacerdotes 15:11 And David called for Zadok and Abiathar the

priests, and for the Levites, for Uriel, Asaiah, and Joel, Shemaiah, and Eliel, and Amminadab,

et Leuitas Vrihel Asaiam Iohel Semeiam Elihel et Aminadab 15:12 et dixit ad eos uos qui estis principes familiarum leui- 15:12 And said unto them, Ye are the chief of the fa-

thers of the Levites: sanctify yourselves, both ye and your brethren, that ye may bring up the ark of the LORD God

ticarum sanctificamini cum fratribus uestris et adferte arcam of Israel unto the place that I have prepared for it. Domini Dei Israhel ad locum qui ei praeparatus est 15:13 15:13 For because ye did it not at the first, the LORD our

God made a breach upon us, for that we sought him not after the due order.

ne ut a principio quia non eratis praesentes percussit nos Dominus sic et nunc fiat inlicitum quid nobis agentibus 15:14 15:14 So the priests and the Levites sanctified themselves

to bring up the ark of the LORD God of Israel.

sanctificati sunt ergo sacerdotes et Leuitae ut portarent arcam Domini Dei Israhel 15:15 et tulerunt filii Leui arcam 15:15 And the children of the Levites bare the ark of God

upon their shoulders with the staves thereon, as Moses commanded according to the word of the LORD.

Dei sicut praeceperat Moses iuxta uerbum Domini umeris suis in uectibus 15:16 dixit quoque Dauid principibus Le- 15:16 And David spake to the chief of the Levites to ap-

point their brethren to be the singers with instruments of musick, psalteries and harps and cymbals, sounding, by

uitarum ut constituerent de fratribus suis cantores in orga- lifting up the voice with joy. nis musicorum nablis uidelicet et lyris et cymbalis ut resonaret in excelsum sonitus laetitiae 15:17 constitueruntque 15:17 So the Levites appointed Heman the son of Joel; and

of his brethren, Asaph the son of Berechiah; and of the sons of Merari their brethren, Ethan the son of Kushaiah;

Leuitas Heman filium Iohel et de fratribus eius Asaph filium Barachiae de filiis uero Merari fratribus eorum Ethan

filium Casaiae 15:18 et cum eis fratres eorum in secundo 15:18 And with them their brethren of the second degree, Zechariah, Ben, and Jaaziel, and Shemiramoth, and Jehiel, and Unni, Eliab, and Benaiah, and Maaseiah, and Mattithiah, and Elipheleh, and Mikneiah, and Obededom, and Jeiel, the porters.

ordine Zacchariam et Ben et Iazihel et Semiramoth et Iahihel et Ani Eliab et Banaiam et Maasiam et Matthathiam et Eliphalu et Macheniam et Obededom et Ieihel ianitores 15:19 porro cantores Heman Asaph et Ethan in cymbalis ae- 15:19 So the singers, Heman, Asaph, and Ethan, were ap-

pointed to sound with cymbals of brass;

neis concrepantes 15:20 Zaccharias autem et Ozihel et Se- 15:20 And Zechariah, and Aziel, and Shemiramoth, and

Jehiel, and Unni, and Eliab, and Maaseiah, and Benaiah, with psalteries on Alamoth;

miramoth et Iahihel et Ani et Eliab et Maasias et Banaias in nablis arcana cantabant 15:21 porro Matthathias et Eliphalu 15:21 And Mattithiah, and Elipheleh, and Mikneiah, and

Obededom, and Jeiel, and Azaziah, with harps on the Sheminith to excel.

et Machenias et Obededom et Ieihel et Ozaziu in citharis pro octaua canebant epinikion 15:22 Chonenias autem princeps 15:22 And Chenaniah, chief of the Levites, was for song:

he instructed about the song, because he was skilful.

Biblia Sacra Vulgata

Leuitarum prophetiae praeerat ad praecinendam melodiam

15:23 And Berechiah and Elkanah were doorkeepers for erat quippe ualde sapiens 15:23 et Barachias et Helcana ia-

the ark.

15:24 And Shebaniah, and Jehoshaphat, and Nethaneel, nitores arcae 15:24 porro Sebenias et Iosaphat et Nathana-

and Amasai, and Zechariah, and Benaiah, and Eliezer, the priests, did blow with the trumpets before the ark of God:

and Obededom and Jehiah were doorkeepers for the ark. hel et Amasai et Zaccharias et Banaias et Eliezer sacerdotes

clangebant tubis coram arca Dei et Obededom et Ahias erant

15:25 So David, and the elders of Israel, and the captains ianitores arcae 15:25 igitur Dauid et maiores natu Israhel

over thousands, went to bring up the ark of the covenant of the LORD out of the house of Obededom with joy.

et tribuni ierunt ad deportandam arcam foederis Domini de

15:26 And it came to pass, when God helped the Levites domo Obededom cum laetitia 15:26 cumque adiuuisset De-

that bare the ark of the covenant of the LORD, that they offered seven bullocks and seven rams.

us Leuitas qui portabant arcam foederis Domini immolaban-

15:27 And David was clothed with a robe of fine linen, tur septem tauri et septem arietes 15:27 porro Dauid erat in-

and all the Levites that bare the ark, and the singers, and Chenaniah the master of the song with the singers: David

also had upon him an ephod of linen. dutus stola byssina et uniuersi Leuitae qui portabant arcam

cantoresque et Chonenias princeps prophetiae inter cantores

15:28 Thus all Israel brought up the ark of the covenant Dauid autem indutus erat etiam ephod lineo 15:28 uniuer-

of the LORD with shouting, and with sound of the cornet, and with trumpets, and with cymbals, making a noise with

psalteries and harps. susque Israhel deducebant arcam foederis Domini in iubilo

et sonitu bucinae et tubis et cymbalis et nablis et citharis

15:29 And it came to pass, as the ark of the covenant of the concrepantes 15:29 cumque peruenisset arca foederis Do-

LORD came to the city of David, that Michal, the daughter of Saul looking out at a window saw king David dancing

and playing: and she despised him in her heart. mini usque ad ciuitatem Dauid Michol filia Saul prospiciens per fenestram uidit regem Dauid saltantem atque ludentem et despexit eum in corde suo

16:1 So they brought the ark of God, and set it in the midst 16:1 adtulerunt igitur arcam Dei et constituerunt eam in

of the tent that David had pitched for it: and they offered burnt sacrifices and peace offerings before God.

medio tabernaculi quod tetenderat ei Dauid et obtulerunt ho-

16:2 And when David had made an end of offering the locausta et pacifica coram Deo 16:2 cumque conplesset Da-

burnt offerings and the peace offerings, he blessed the people in the name of the LORD.

uid offerens holocausta et pacifica benedixit populo in no-

16:3 And he dealt to every one of Israel, both man and mine Domini 16:3 et diuisit uniuersis per singulos a uiro

woman, to every one a loaf of bread, and a good piece of flesh, and a flagon of wine.

usque ad mulierem tortam panis et partem assae carnis bu-

16:4 And he appointed certain of the Levites to minister bulae et frixam oleo similam 16:4 constituitque coram arca

before the ark of the LORD, and to record, and to thank and praise the LORD God of Israel:

Domini de Leuitis qui ministrarent et recordarentur operum eius et glorificarent atque laudarent Dominum Deum Israhel

16:5 Asaph the chief, and next to him Zechariah, Jeiel, 16:5 Asaph principem et secundum eius Zacchariam porro

and Shemiramoth, and Jehiel, and Mattithiah, and Eliab, and Benaiah, and Obededom: and Jeiel with psalteries and

with harps; but Asaph made a sound with cymbals; Iahihel et Semiramoth et Ieihel et Matthathiam et Eliab et

Banaiam et Obededom et Ieihel super organa psalterii et ly-

16:6 Benaiah also and Jahaziel the priests with trumpets ras Asaph autem ut cymbalis personaret 16:6 Banaiam uero continually before the ark of the covenant of God.

et Azihel sacerdotes canere tuba iugiter coram arca foederis Domini 16:7 in illo die fecit Dauid principem ad con- 16:7 Then on that day David delivered first this psalm to

thank the LORD into the hand of Asaph and his brethren.

fitendum Domino Asaph et fratres eius 16:8 confitemini 16:8 Give thanks unto the LORD, call upon his name,

make known his deeds among the people.

Domino inuocate nomen eius notas facite in populis adinuentiones illius 16:9 canite ei et psallite et narrate omnia 16:9 Sing unto him, sing psalms unto him, talk ye of all

his wondrous works.

mirabilia eius 16:10 laudate nomen sanctum eius laetetur 16:10 Glory ye in his holy name: let the heart of them

rejoice that seek the LORD.

cor quaerentium Dominum 16:11 quaerite Dominum et uir- 16:11 Seek the LORD and his strength, seek his face con-

tinually.

tutem eius quaerite faciem eius semper 16:12 recordamini 16:12 Remember his marvellous works that he hath done,

his wonders, and the judgments of his mouth;

mirabilium eius quae fecit signorum illius et iudiciorum oris eius 16:13 semen Israhel serui eius filii Iacob electi illius 16:13 O ye seed of Israel his servant, ye children of Jacob,

his chosen ones.

16:14 ipse Dominus Deus noster in uniuersa terra iudicia 16:14 He is the LORD our God; his judgments are in all

the earth.

eius 16:15 recordamini in sempiternum pacti eius sermonis 16:15 Be ye mindful always of his covenant; the word

which he commanded to a thousand generations;

quem praecepit in mille generationes 16:16 quem pepigit 16:16 Even of the covenant which he made with Abraham,

and of his oath unto Isaac;

cum Abraham et iuramenti illius cum Isaac 16:17 et con- 16:17 And hath confirmed the same to Jacob for a law, and

to Israel for an everlasting covenant,

stituit illud Iacob in praeceptum et Israhel in pactum sempiternum 16:18 dicens tibi dabo terram Chanaan funiculum 16:18 Saying, Unto thee will I give the land of Canaan, the

lot of your inheritance;

hereditatis uestrae 16:19 cum essent pauci numero parui et 16:19 When ye were but few, even a few, and strangers in

it.

coloni eius 16:20 et transierunt de gente in gentem et de re- 16:20 And when they went from nation to nation, and from

one kingdom to another people;

gno ad populum alterum 16:21 non dimisit quemquam ca- 16:21 He suffered no man to do them wrong: yea, he re-

proved kings for their sakes,

lumniari eos sed increpuit pro eis reges 16:22 nolite tangere 16:22 Saying, Touch not mine anointed, and do my

prophets no harm.

christos meos et in prophetis meis nolite malignari 16:23 16:23 Sing unto the LORD, all the earth; shew forth from

day to day his salvation.

canite Domino omnis terra adnuntiate ex die in diem salutare eius 16:24 narrate in gentibus gloriam eius in cunctis popu- 16:24 Declare his glory among the heathen; his marvel-

lous works among all nations.

lis mirabilia illius 16:25 quia magnus Dominus et laudabilis 16:25 For great is the LORD, and greatly to be praised: he

also is to be feared above all gods.

nimis et horribilis super omnes deos 16:26 omnes enim dii 16:26 For all the gods of the people are idols: but the

LORD made the heavens.

populorum idola Dominus autem caelos fecit 16:27 confes- 16:27 Glory and honour are in his presence; strength and

gladness are in his place.

sio et magnificentia coram eo fortitudo et gaudium in loco eius 16:28 adferte Domino familiae populorum adferte Do- 16:28 Give unto the LORD, ye kindreds of the people, give

unto the LORD glory and strength.

mino gloriam et imperium 16:29 date Domino gloriam no- 16:29 Give unto the LORD the glory due unto his name:

bring an offering, and come before him: worship the LORD in the beauty of holiness.

mini eius leuate sacrificium et uenite in conspectu eius et

adorate Dominum in decore sancto 16:30 commoueatur a 16:30 Fear before him, all the earth: the world also shall

be stable, that it be not moved.

facie illius omnis terra ipse enim fundauit orbem inmobilem 16:31 laetentur caeli et exultet terra et dicant in natio- 16:31 Let the heavens be glad, and let the earth rejoice:

and let men say among the nations, The LORD reigneth.

nibus Dominus regnauit 16:32 tonet mare et plenitudo eius 16:32 Let the sea roar, and the fulness thereof: let the fields rejoice, and all that is therein.

Biblia Sacra Vulgata

16:33 Then shall the trees of the wood sing out at the pres- exultent agri et omnia quae in eis sunt 16:33 tunc laudaence of the LORD, because he cometh to judge the earth.

bunt ligna saltus coram Domino quia uenit iudicare terram

16:34 O give thanks unto the LORD; for he is good; for 16:34 confitemini Domino quoniam bonus quoniam in aeter-

his mercy endureth for ever.

16:35 And say ye, Save us, O God of our salvation, and num misericordia eius 16:35 et dicite salua nos Deus sal-

gather us together, and deliver us from the heathen, that we may give thanks to thy holy name, and glory in thy

praise. uator noster et congrega nos et erue de gentibus ut confi-

teamur nomini sancto tuo et exultemus in carminibus tuis

16:36 Blessed be the LORD God of Israel for ever and 16:36 benedictus Dominus Deus Israhel ab aeterno usque in

ever. And all the people said, Amen, and praised the LORD.

aeternum et dicat omnis populus amen et hymnus Domino

16:37 So he left there before the ark of the covenant of the 16:37 dereliquit itaque ibi coram arca foederis Domini Asaph

LORD Asaph and his brethren, to minister before the ark continually, as every day’s work required:

et fratres eius ut ministrarent in conspectu arcae iugiter per

16:38 And Obededom with their brethren, threescore and singulos dies et uices suas 16:38 porro Obededom et frat-

eight; Obededom also the son of Jeduthun and Hosah to be porters:

res eius sexaginta octo et Obededom filium Idithun et Osa

16:39 And Zadok the priest, and his brethren the priests, constituit ianitores 16:39 Sadoc autem sacerdotem et frat-

before the tabernacle of the LORD in the high place that was at Gibeon,

res illius sacerdotes coram tabernaculo Domini in excelso

16:40 To offer burnt offerings unto the LORD upon the al- quod erat in Gabaon 16:40 ut offerrent holocausta Domino

tar of the burnt offering continually morning and evening, and to do according to all that is written in the law of the

LORD, which he commanded Israel; super altare holocaustomatis iugiter mane et uespere iuxta

omnia quae scripta sunt in lege Domini quam praecepit Isra-

16:41 And with them Heman and Jeduthun, and the rest heli 16:41 et post eum Heman et Idithun et reliquos electos

that were chosen, who were expressed by name, to give thanks to the LORD, because his mercy endureth for ever;

unumquemque uocabulo suo ad confitendum Domino quo-

16:42 And with them Heman and Jeduthun with trumpets niam in aeternum misericordia eius 16:42 Heman quoque

and cymbals for those that should make a sound, and with musical instruments of God. And the sons of Jeduthun

were porters. et Idithun canentes tuba et quatientes cymbala et omnia mu-

sicorum organa ad canendum Deo filios autem Idithun fecit

16:43 And all the people departed every man to his house: esse portarios 16:43 reuersusque est omnis populus in do-

and David returned to bless his house.

mum suam et Dauid ut benediceret etiam domui suae

17:1 Now it came to pass, as David sat in his house, that 17:1 cum autem habitaret Dauid in domo sua dixit ad Nat-

David said to Nathan the prophet, Lo, I dwell in an house of cedars, but the ark of the covenant of the LORD re-

maineth under curtains. han prophetam ecce ego habito in domo cedrina arca autem 17:2 Then Nathan said unto David, Do all that is in thine foederis Domini sub pellibus est 17:2 et ait Nathan ad Da-

heart; for God is with thee.

uid omnia quae in corde tuo sunt fac Deus enim tecum est

17:3 And it came to pass the same night, that the word of 17:3 igitur nocte illa factus est sermo Dei ad Nathan dicens

God came to Nathan, saying,

17:4 Go and tell David my servant, Thus saith the LORD, 17:4 uade et loquere Dauid seruo meo haec dicit Dominus

Thou shalt not build me an house to dwell in: 17:5 For I have not dwelt in an house since the day that I non aedificabis tu mihi domum ad habitandum 17:5 neque brought up Israel unto this day; but have gone from tent to tent, and from one tabernacle to another.

enim mansi in domo ex eo tempore quo eduxi Israhel usque ad hanc diem sed fui semper mutans loca tabernaculi et

Liber I Paralipomenon

in tentorio 17:6 manens cum omni Israhel numquid locutus 17:6 Wheresoever I have walked with all Israel, spake I a

word to any of the judges of Israel, whom I commanded to feed my people, saying, Why have ye not built me an

sum saltim uni iudicum Israhel quibus praeceperam ut pa- house of cedars? scerent populum meum et dixi quare non aedificastis mihi domum cedrinam 17:7 nunc itaque sic loqueris ad seruum 17:7 Now therefore thus shalt thou say unto my servant

David, Thus saith the LORD of hosts, I took thee from the sheepcote, even from following the sheep, that thou

meum Dauid haec dicit Dominus exercituum ego tuli te cum shouldest be ruler over my people Israel: in pascuis sequereris gregem ut esses dux populi mei Israhel 17:8 et fui tecum quocumque perrexisti et interfeci om- 17:8 And I have been with thee whithersoever thou hast

walked, and have cut off all thine enemies from before thee, and have made thee a name like the name of the great

nes inimicos tuos coram te fecique tibi nomen quasi unius men that are in the earth. magnorum qui celebrantur in terra 17:9 et dedi locum po- 17:9 Also I will ordain a place for my people Israel, and

will plant them, and they shall dwell in their place, and shall be moved no more; neither shall the children of

pulo meo Israhel plantabitur et habitabit in eo et ultra non wickedness waste them any more, as at the beginning, commouebitur nec filii iniquitatis adterent eos sicut a principio 17:10 ex diebus quibus dedi iudices populo meo Isra- 17:10 And since the time that I commanded judges to be

over my people Israel. Moreover I will subdue all thine enemies. Furthermore I tell thee that the LORD will build

hel et humiliaui uniuersos inimicos tuos adnuntio ergo tibi thee an house. quod aedificaturus sit domum tibi Dominus 17:11 cumque 17:11 And it shall come to pass, when thy days be expired

that thou must go to be with thy fathers, that I will raise up thy seed after thee, which shall be of thy sons; and I will

impleueris dies tuos ut uadas ad patres tuos suscitabo semen establish his kingdom. tuum post te quod erit de filiis tuis et stabiliam regnum eius 17:12 ipse aedificabit mihi domum et firmabo solium eius us- 17:12 He shall build me an house, and I will stablish his

throne for ever.

que in aeternum 17:13 ego ero ei in patrem et ipse erit mihi 17:13 I will be his father, and he shall be my son: and I

will not take my mercy away from him, as I took it from him that was before thee:

in filium et misericordiam meam non auferam ab eo sicut abstuli ab eo qui ante te fuit 17:14 et statuam eum in domo 17:14 But I will settle him in mine house and in my king-

dom for ever: and his throne shall be established for evermore.

mea et in regno meo usque in sempiternum et thronus eius erit firmissimus in perpetuum 17:15 iuxta omnia uerba haec 17:15 According to all these words, and according to all

this vision, so did Nathan speak unto David.

et iuxta uniuersam uisionem istam sic locutus est Nathan ad Dauid 17:16 cumque uenisset rex Dauid et sedisset coram 17:16 And David the king came and sat before the LORD,

and said, Who am I, O LORD God, and what is mine house, that thou hast brought me hitherto?

Domino dixit quis ego sum Domine Deus et quae domus mea ut praestares mihi talia 17:17 sed et hoc parum uisum 17:17 And yet this was a small thing in thine eyes, O God;

for thou hast also spoken of thy servant’s house for a great while to come, and hast regarded me according to the es-

est in conspectu tuo ideoque locutus es super domum serui tate of a man of high degree, O LORD God. tui etiam in futurum et fecisti me spectabilem super omnes homines Domine Deus meus 17:18 quid ultra addere potest 17:18 What can David speak more to thee for the honour

of thy servant? for thou knowest thy servant.

Dauid cum ita glorificaueris seruum tuum et cognoueris eum 17:19 Domine propter famulum tuum iuxta cor tuum fecisti 17:19 O LORD, for thy servant’s sake, and according to

thine own heart, hast thou done all this greatness, in making known all these great things.

omnem magnificentiam hanc et nota esse uoluisti uniuersa magnalia 17:20 Domine non est similis tui et non est alius 17:20 O LORD, there is none like thee, neither is there any

God beside thee, according to all that we have heard with our ears.

Biblia Sacra Vulgata

deus absque te ex omnibus quos audiuimus auribus nostris

17:21 And what one nation in the earth is like thy people 17:21 quis autem est alius ut populus tuus Israhel gens una in

Israel, whom God went to redeem to be his own people, to make thee a name of greatness and terribleness, by driving out nations from before thy people whom thou hast redeemed out of Egypt?

terra ad quam perrexit Deus ut liberaret et faceret populum sibi et magnitudine sua atque terroribus eiceret nationes a

17:22 For thy people Israel didst thou make thine own peo- facie eius quem de Aegypto liberarat 17:22 et posuisti po-

ple for ever; and thou, LORD, becamest their God.

pulum tuum Israhel tibi in populum usque in aeternum et

17:23 Therefore now, LORD, let the thing that thou hast tu Domine factus es Deus eius 17:23 nunc igitur Domine spoken concerning thy servant and concerning his house be established for ever, and do as thou hast said.

sermo quem locutus es famulo tuo et super domum eius con-

17:24 Let it even be established, that thy name may be firmetur in perpetuum et fac sicut locutus es 17:24 perma-

magnified for ever, saying, The LORD of hosts is the God of Israel, even a God to Israel: and let the house of David

thy servant be established before thee. neatque et magnificetur nomen tuum usque in sempiternum

et dicatur Dominus exercituum Deus Israhel et domus Dauid

17:25 For thou, O my God, hast told thy servant that thou serui eius permanens coram eo 17:25 tu enim Domine Deus wilt build him an house: therefore thy servant hath found in his heart to pray before thee.

meus reuelasti auriculam serui tui ut aedificares ei domum et

17:26 And now, LORD, thou art God, and hast promised idcirco inuenit seruus tuus fiduciam ut oret coram te 17:26

this goodness unto thy servant:

nunc ergo Domine tu es Deus et locutus es ad seruum tuum

17:27 Now therefore let it please thee to bless the house tanta beneficia 17:27 et coepisti benedicere domui serui tui

of thy servant, that it may be before thee for ever: for thou blessest, O LORD, and it shall be blessed for ever.

ut sit semper coram te te enim Domine benedicente benedicta erit in perpetuum

18:1 Now after this it came to pass, that David smote 18:1 factum est autem post haec ut percuteret Dauid Phi-

the Philistines, and subdued them, and took Gath and her towns out of the hand of the Philistines.

listhim et humiliaret eos et tolleret Geth et filias eius de

18:2 And he smote Moab; and the Moabites became manu Philisthim 18:2 percuteretque Moab et fierent Mo-

David’s servants, and brought gifts.

18:3 And David smote Hadarezer king of Zobah unto abitae serui Dauid offerentes ei munera 18:3 eo tempore

Hamath, as he went to stablish his dominion by the river Euphrates.

percussit Dauid etiam Adadezer regem Suba regionis Emath quando perrexit ut dilataret imperium suum usque ad flumen

18:4 And David took from him a thousand chariots, and Eufraten 18:4 cepit ergo Dauid mille quadrigas eius et se-

seven thousand horsemen, and twenty thousand footmen: David also houghed all the chariot horses, but reserved of

them an hundred chariots. ptem milia equites ac uiginti milia uirorum peditum subner-

uauitque omnes equos curruum exceptis centum quadrigis

18:5 And when the Syrians of Damascus came to help quas reseruauit sibi 18:5 superuenit autem et Syrus damas-

Hadarezer king of Zobah, David slew of the Syrians two and twenty thousand men.

cenus ut auxilium praeberet Adadezer regi Suba sed et huius

18:6 Then David put garrisons in Syriadamascus; and the percussit Dauid uiginti duo milia uirorum 18:6 et posuit mi-

Syrians became David’s servants, and brought gifts. Thus the LORD preserved David whithersoever he went.

lites in Damasco ut Syria quoque seruiret sibi et offerret munera adiuuitque eum Dominus in cunctis ad quae perrexerat

18:7 And David took the shields of gold that were on the 18:7 tulit quoque Dauid faretras aureas quas habuerant serui

servants of Hadarezer, and brought them to Jerusalem.

Liber I Paralipomenon

Adadezer et adtulit eas in Hierusalem 18:8 necnon de The- 18:8 Likewise from Tibhath, and from Chun, cities of

Hadarezer, brought David very much brass, wherewith Solomon made the brasen sea, and the pillars, and the ves-

bath et Chun urbibus Adadezer aeris plurimum de quo fe- sels of brass. cit Salomon mare aeneum et columnas et uasa aenea 18:9 18:9 Now when Tou king of Hamath heard how David had

smitten all the host of Hadarezer king of Zobah;

quod cum audisset Thou rex Emath percussisse uidelicet Dauid omnem exercitum Adadezer regis Suba 18:10 misit Ad- 18:10 He sent Hadoram his son to king David, to enquire

of his welfare, and to congratulate him, because he had fought against Hadarezer, and smitten him; (for Hadarezer had war with Tou;) and with him all manner of vessels of gold and silver and brass.

uram filium suum ad regem Dauid ut postularet ab eo pacem et congratularetur ei eo quod expugnasset et percussisset Adadezer aduersarius quippe Thou erat Adadezer 18:11 sed et 18:11 Them also king David dedicated unto the LORD,

with the silver and the gold that he brought from all these nations; from Edom, and from Moab, and from the children of Ammon, and from the Philistines, and from Amalek.

omnia uasa aurea et argentea et aenea consecrauit rex Dauid Domino cum argento et auro quod tulerat ex uniuersis gentibus tam de Idumea et Moab et filiis Ammon quam de Philisthim et Amalech 18:12 Abisai uero filius Saruiae percussit 18:12 Moreover Abishai the son of Zeruiah slew of the

Edomites in the valley of salt eighteen thousand.

Edom in ualle Salinarum decem et octo milia 18:13 et con- 18:13 And he put garrisons in Edom; and all the Edomites

became David’s servants. Thus the LORD preserved David whithersoever he went.

stituit in Edom praesidium ut seruiret Idumea Dauid salua-

uitque Dominus Dauid in cunctis ad quae perrexerat 18:14 18:14 So David reigned over all Israel, and executed judg-

ment and justice among all his people.

regnauit ergo Dauid super uniuersum Israhel et faciebat iudicium atque iustitiam cuncto populo suo 18:15 porro Ioab 18:15 And Joab the son of Zeruiah was over the host; and

Jehoshaphat the son of Ahilud, recorder.

filius Saruiae erat super exercitum et Iosaphat filius Ahilud a commentariis 18:16 Sadoc autem filius Ahitob et Ahime- 18:16 And Zadok the son of Ahitub, and Abimelech the

son of Abiathar, were the priests; and Shavsha was scribe;

lech filius Abiathar sacerdotes et Susa scriba 18:17 Banaias 18:17 And Benaiah the son of Jehoiada was over the

Cherethites and the Pelethites; and the sons of David were chief about the king.

uero filius Ioiada super legiones Cherethi et Felethi porro filii Dauid primi ad manum regis

19:1 accidit autem ut moreretur Naas rex filiorum Ammon 19:1 Now it came to pass after this, that Nahash the king

of the children of Ammon died, and his son reigned in his stead.

et regnaret filius eius pro eo 19:2 dixitque Dauid faciam mi-

19:2 And David said, I will shew kindness unto Hanun the son of Nahash, because his father shewed kindness to me. And David sent messengers to comfort him concerning his father. So the servants of David came into the land of the children of Ammon to Hanun, to comfort him.

sericordiam cum Hanon filio Naas praestitit enim pater eius mihi gratiam misitque Dauid nuntios ad consolandum eum super morte patris sui qui cum peruenissent in terram filiorum Ammon ut consolarentur Hanon 19:3 dixerunt princi- 19:3 But the princes of the children of Ammon said to

Hanun, Thinkest thou that David doth honour thy father, that he hath sent comforters unto thee? are not his servants come unto thee for to search, and to overthrow, and to spy out the land?

pes filiorum Ammon ad Hanon tu forsitan putas quod Dauid honoris causa in patrem tuum miserit qui consolentur te nec animaduertis quod ut explorent et inuestigent et scrutentur terram tuam uenerint ad te serui eius 19:4 igitur 19:4 Wherefore Hanun took David’s servants, and shaved

them, and cut off their garments in the midst hard by their buttocks, and sent them away.

Hanon pueros Dauid decaluauit et rasit et praecidit tunicas

Biblia Sacra Vulgata

19:5 Then there went certain, and told David how the men eorum a natibus usque ad pedes et dimisit eos 19:5 qui cum

were served. And he sent to meet them: for the men were greatly ashamed. And the king said, Tarry at Jericho until

your beards be grown, and then return. abissent et hoc mandassent Dauid misit in occursum eorum grandem enim contumeliam sustinuerant et praecepit ut manerent in Hiericho donec cresceret barba eorum et tunc re-

19:6 And when the children of Ammon saw that they had uerterentur 19:6 uidentes autem filii Ammon quod iniuriam

made themselves odious to David, Hanun and the children of Ammon sent a thousand talents of silver to hire them chariots and horsemen out of Mesopotamia, and out of Syriamaachah, and out of Zobah.

fecissent Dauid tam Hanon quam reliquus populus miserunt mille talenta argenti ut conducerent sibi de Mesopotamia et

19:7 So they hired thirty and two thousand chariots, and de Syria Macha et de Suba currus et equites 19:7 conduxethe king of Maachah and his people; who came and pitched before Medeba. And the children of Ammon gathered themselves together from their cities, and came to battle.

runtque triginta duo milia curruum et regem Macha cum populo eius qui cum uenissent castrametati sunt e regione Medaba filii quoque Ammon congregati de urbibus suis uene-

19:8 And when David heard of it, he sent Joab, and all the runt ad bellum 19:8 quod cum audisset Dauid misit Ioab host of the mighty men.

19:9 And the children of Ammon came out, and put the et omnem exercitum uirorum fortium 19:9 egressique filii

battle in array before the gate of the city: and the kings that were come were by themselves in the field.

Ammon direxerunt aciem iuxta portam ciuitatis reges autem

19:10 Now when Joab saw that the battle was set against qui ad auxilium uenerant separatim in agro steterunt 19:10

him before and behind, he chose out of all the choice of Israel, and put them in array against the Syrians.

igitur Ioab intellegens bellum et ex aduerso et post tergum contra se fieri elegit uiros fortissimos de uniuerso Israhel et

19:11 And the rest of the people he delivered unto the hand perrexit contra Syrum 19:11 reliquam autem partem populi

of Abishai his brother, and they set themselves in array against the children of Ammon.

dedit sub manu Abisai fratris sui et perrexerunt contra filios

19:12 And he said, If the Syrians be too strong for me, Ammon 19:12 dixitque si uicerit me Syrus auxilio eris mihi

then thou shalt help me: but if the children of Ammon be too strong for thee, then I will help thee.

sin autem superauerint te filii Ammon ero tibi in praesidium

19:13 Be of good courage, and let us behave ourselves 19:13 confortare et agamus uiriliter pro populo nostro et pro

valiantly for our people, and for the cities of our God: and let the LORD do that which is good in his sight.

urbibus Dei nostri Dominus autem quod in conspectu suo

19:14 So Joab and the people that were with him drew nigh bonum est faciet 19:14 perrexit ergo Ioab et populus qui

before the Syrians unto the battle; and they fled before him.

cum eo erat contra Syrum ad proelium et fugauit eos 19:15

19:15 And when the children of Ammon saw that the Syrians were fled, they likewise fled before Abishai his brother, and entered into the city. Then Joab came to Jerusalem.

porro filii Ammon uidentes quod fugisset Syrus ipsi quoque fugerunt Abisai fratrem eius et ingressi sunt ciuitatem reuer-

19:16 And when the Syrians saw that they were put to the susque est etiam Ioab in Hierusalem 19:16 uidens autem

worse before Israel, they sent messengers, and drew forth the Syrians that were beyond the river: and Shophach the

captain of the host of Hadarezer went before them. Syrus quod cecidisset coram Israhel misit nuntios et adduxit

Syrum qui erat trans Fluuium Sophach autem princeps mi-

19:17 And it was told David; and he gathered all Israel, litiae Adadezer erat dux eorum 19:17 quod cum nuntiatum

and passed over Jordan, and came upon them, and set the battle in array against them. So when David had put the

battle in array against the Syrians, they fought with him. esset Dauid congregauit uniuersum Israhel et transiuit Iordanem inruitque in eos et direxit ex aduerso aciem illis con-

Liber I Paralipomenon

tra pugnantibus 19:18 fugit autem Syrus Israhel et interfecit 19:18 But the Syrians fled before Israel; and David slew

of the Syrians seven thousand men which fought in chariots, and forty thousand footmen, and killed Shophach the

Dauid de Syris septem milia curruum et quadraginta milia captain of the host. peditum et Sophach exercitus principem 19:19 uidentes au- 19:19 And when the servants of Hadarezer saw that they

were put to the worse before Israel, they made peace with David, and became his servants: neither would the Syrians

tem serui Adadezer se ab Israhel esse superatos transfuge- help the children of Ammon any more. runt ad Dauid et seruierunt ei noluitque ultra Syria auxilium praebere filiis Ammon

20:1 factum est autem post anni circulum eo tempore quo 20:1 And it came to pass, that after the year was expired, at the time that kings go out to battle, Joab led forth the power of the army, and wasted the country of the children of Ammon, and came and besieged Rabbah. But David tarried at Jerusalem. And Joab smote Rabbah, and destroyed it.

solent reges ad bella procedere congregauit Ioab exercitum et robur militiae et uastauit terram filiorum Ammon perrexitque et obsedit Rabba porro Dauid manebat in Hierusalem quando Ioab percussit Rabba et destruxit eam 20:2 tulit au- 20:2 And David took the crown of their king from off his

head, and found it to weigh a talent of gold, and there were precious stones in it; and it was set upon David’s head: and

tem Dauid coronam Melchom de capite eius et inuenit in ea he brought also exceeding much spoil out of the city. auri pondo talentum et pretiosissimas gemmas fecitque sibi inde diadema manubias quoque urbis plurimas tulit 20:3 20:3 And he brought out the people that were in it, and

cut them with saws, and with harrows of iron, and with axes. Even so dealt David with all the cities of the children of Ammon. And David and all the people returned to Jerusalem.

populum autem qui erat in ea eduxit et fecit super eos tribulas et trahas et ferrata carpenta transire ita ut dissicarentur et contererentur sic fecit Dauid cunctis urbibus filiorum Ammon et reuersus est cum omni populo suo in Hierusalem 20:4 post haec initum est bellum in Gazer aduersus Philist- 20:4 And it came to pass after this, that there arose war

at Gezer with the Philistines; at which time Sibbechai the Hushathite slew Sippai, that was of the children of the gi-

heos in quo percussit Sobbochai Vsathites Saphai de genere ant: and they were subdued. Raphaim et humiliauit eos 20:5 aliud quoque bellum ge- 20:5 And there was war again with the Philistines; and

Elhanan the son of Jair slew Lahmi the brother of Goliath the Gittite, whose spear staff was like a weaver’s beam.

stum est aduersum Philistheos in quo percussit Adeodatus filius Saltus Lehemites fratrem Goliath Getthei cuius hastae lignum erat quasi liciatorium texentium 20:6 sed et aliud 20:6 And yet again there was war at Gath, where was a

man of great stature, whose fingers and toes were four and twenty, six on each hand, and six on each foot and he also

bellum accidit in Geth in quo fuit homo longissimus ha- was the son of the giant. bens digitos senos id est simul uiginti quattuor qui et ipse de Rapha fuerat stirpe generatus 20:7 hic blasphemauit Is- 20:7 But when he defied Israel, Jonathan the son of

Shimea David’s brother slew him. These were born unto the giant in Gath; and they fell by the hand of David, and

rahel et percussit eum Ionathan filius Sammaa fratris Dauid by the hand of his servants. hii sunt filii Rapha in Geth qui ceciderunt in manu Dauid et seruorum eius

21:1 consurrexit autem Satan contra Israhel et incitauit Da- 21:1 And Satan stood up against Israel, and provoked

David to number Israel.

uid ut numeraret Israhel 21:2 dixitque Dauid ad Ioab et ad 21:2 And David said to Joab and to the rulers of the people,

Go, number Israel from Beersheba even to Dan; and bring the number of them to me, that I may know it.

principes populi ite et numerate Israhel a Bersabee usque

Biblia Sacra Vulgata

21:3 And Joab answered, The LORD make his people an Dan et adferte mihi numerum ut sciam 21:3 responditque

hundred times so many more as they be: but, my lord the king, are they not all my lord’s servants? why then doth my lord require this thing? why will he be a cause of trespass to Israel?

Ioab augeat Dominus populum suum centuplum quam sunt nonne domine mi rex omnes serui tui sunt quare hoc quaerit

21:4 Nevertheless the king’s word prevailed against Joab. dominus meus quod in peccatum reputetur Israheli 21:4 sed Wherefore Joab departed, and went throughout all Israel, and came to Jerusalem.

sermo regis magis praeualuit egressusque est Ioab et circui-

21:5 And Joab gave the sum of the number of the people uit uniuersum Israhel et reuersus est Hierusalem 21:5 de-

unto David. And all they of Israel were a thousand thousand and an hundred thousand men that drew sword: and Judah was four hundred threescore and ten thousand men that drew sword.

ditque Dauid numerum eorum quos circumierat et inuentus est omnis Israhel numerus mille milia et centum milia uirorum educentium gladium de Iuda autem trecenta septuaginta

21:6 But Levi and Benjamin counted he not among them: milia bellatorum 21:6 nam Leui et Beniamin non numera-

for the king’s word was abominable to Joab.

21:7 And God was displeased with this thing; therefore he uit eo quod inuitus exsequeretur regis imperium 21:7 disp-

smote Israel.

21:8 And David said unto God, I have sinned greatly, be- licuit autem Deo quod iussum erat et percussit Israhel 21:8

cause I have done this thing: but now, I beseech thee, do away the iniquity of thy servant; for I have done very fool-

ishly. dixitque Dauid ad Deum peccaui nimis ut hoc facerem ob-21:9 And the LORD spake unto Gad, David’s seer, saying, secro aufer iniquitatem serui tui quia insipienter egi 21:9 et 21:10 Go and tell David, saying, Thus saith the LORD, locutus est Dominus ad Gad uidentem Dauid dicens 21:10

I offer thee three things: choose thee one of them, that I may do it unto thee.

uade et loquere ad Dauid et dic haec dicit Dominus trium tibi

21:11 So Gad came to David, and said unto him, Thus optionem do unum quod uolueris elige et faciam tibi 21:11

saith the LORD, Choose thee

cumque uenisset Gad ad Dauid dixit ei haec dicit Dominus

21:12 Either three years’ famine; or three months to be elige quod uolueris 21:12 aut tribus annis pestilentiam aut

destroyed before thy foes, while that the sword of thine enemies overtaketh thee; or else three days the sword of the LORD, even the pestilence, in the land, and the angel of the LORD destroying throughout all the coasts of Israel. Now therefore advise thyself what word I shall bring again to him that sent me.

tribus mensibus fugere te hostes tuos et gladium eorum non posse euadere aut tribus diebus gladium Domini et mortem uersari in terra et angelum Domini interficere in uniuersis finibus Israhel nunc igitur uide quid respondeam ei qui misit

21:13 And David said unto Gad, I am in a great strait: let me 21:13 et dixit Dauid ad Gad ex omni parte me angustiae me fall now into the hand of the LORD; for very great are his mercies: but let me not fall into the hand of man.

premunt sed melius mihi est ut incidam in manus Domini quia multae sunt miserationes eius quam in manus homi-

21:14 So the LORD sent pestilence upon Israel: and there num 21:14 misit ergo Dominus pestilentiam in Israhel et fell of Israel seventy thousand men.

21:15 And God sent an angel unto Jerusalem to destroy ceciderunt de Israhel septuaginta milia uirorum 21:15 misit

it: and as he was destroying, the LORD beheld, and he repented him of the evil, and said to the angel that destroyed, It is enough, stay now thine hand. And the angel of the LORD stood by the threshingfloor of Ornan the Jebusite.

quoque angelum in Hierusalem ut percuteret eam cumque percuteretur uidit Dominus et misertus est super magnitudinem mali et imperauit angelo qui percutiebat sufficit iam cesset manus tua porro angelus Domini stabat iuxta aream

21:16 And David lifted up his eyes, and saw the angel of Ornan Iebusei 21:16 leuansque Dauid oculos suos uidit an-

the LORD stand between the earth and the heaven, having a drawn sword in his hand stretched out over Jerusalem. Then David and the elders of Israel, who were clothed in sackcloth, fell upon their faces.

Liber I Paralipomenon

gelum Domini stantem inter terram et caelum et euaginatum gladium in manu eius et uersum contra Hierusalem et ceciderunt tam ipse quam maiores natu uestiti ciliciis et proni in terram 21:17 dixitque Dauid ad Deum nonne ego sum qui 21:17 And David said unto God, Is it not I that com-

manded the people to be numbered? even I it is that have sinned and done evil indeed; but as for these sheep, what have they done? let thine hand, I pray thee, O LORD my God, be on me, and on my father’s house; but not on thy people, that they should be plagued.

iussi ut numeraretur populus ego qui peccaui ego qui malum feci iste grex quid commeruit Domine Deus meus uertatur obsecro manus tua in me et in domum patris mei populus autem tuus non percutiatur 21:18 angelus autem Domini 21:18 Then the angel of the LORD commanded Gad to say

to David, that David should go up, and set up an altar unto the LORD in the threshingfloor of Ornan the Jebusite.

praecepit Gad ut diceret Dauid et ascenderet extrueretque altare Domino Deo in area Ornan Iebusei 21:19 ascendit ergo 21:19 And David went up at the saying of Gad, which he

spake in the name of the LORD.

Dauid iuxta sermonem Gad quem locutus fuerat ex nomine Domini 21:20 porro Ornan cum suspexisset et uidisset an- 21:20 And Ornan turned back, and saw the angel; and

his four sons with him hid themselves. Now Ornan was threshing wheat.

gelum quattuorque filii eius cum eo absconderunt se nam eo tempore terebat in area triticum 21:21 igitur cum uenisset 21:21 And as David came to Ornan, Ornan looked and

saw David, and went out of the threshingfloor, and bowed himself to David with his face to the ground.

Dauid ad Ornan conspexit eum Ornan et processit ei obuiam de area et adorauit illum pronus in terram 21:22 dixitque ei 21:22 Then David said to Ornan, Grant me the place of

this threshingfloor, that I may build an altar therein unto the LORD: thou shalt grant it me for the full price: that

Dauid da mihi locum areae tuae ut aedificem in ea altare Do- the plague may be stayed from the people. mini ita ut quantum ualet argenti accipias et cesset plaga a populo 21:23 dixit autem Ornan ad Dauid tolle et faciat do- 21:23 And Ornan said unto David, Take it to thee, and

let my lord the king do that which is good in his eyes: lo, I give thee the oxen also for burnt offerings, and the threshing instruments for wood, and the wheat for the meat offering; I give it all.

minus meus rex quodcumque ei placet sed et boues do in holocaustum et tribulas in ligna et triticum in sacrificium omnia libens praebeo 21:24 dixitque ei rex Dauid nequaquam ita 21:24 And king David said to Ornan, Nay; but I will verily

buy it for the full price: for I will not take that which is thine for the LORD, nor offer burnt offerings without cost.

fiet sed argentum dabo quantum ualet neque enim tibi auferre debeo et sic offerre Domino holocausta gratuita 21:25 21:25 So David gave to Ornan for the place six hundred

shekels of gold by weight.

dedit ergo Dauid Ornan pro loco siclos auri iustissimi ponderis sescentos 21:26 et aedificauit ibi altare Domino ob- 21:26 And David built there an altar unto the LORD, and

offered burnt offerings and peace offerings, and called upon the LORD; and he answered him from heaven by

tulitque holocausta et pacifica et inuocauit Dominum et ex- fire upon the altar of burnt offering. audiuit eum in igne de caelo super altare holocausti 21:27 21:27 And the LORD commanded the angel; and he put

up his sword again into the sheath thereof.

praecepitque Dominus angelo et conuertit gladium suum in uaginam 21:28 protinus ergo Dauid uidens quod exaudisset 21:28 At that time when David saw that the LORD had

answered him in the threshingfloor of Ornan the Jebusite, then he sacrificed there.

eum Dominus in area Ornan Iebusei immolauit ibi uictimas 21:29 tabernaculum autem Domini quod fecerat Moses in de- 21:29 For the tabernacle of the LORD, which Moses made

in the wilderness, and the altar of the burnt offering, were at that season in the high place at Gibeon.

serto et altare holocaustorum ea tempestate erat in excelso

Biblia Sacra Vulgata

21:30 But David could not go before it to enquire of God: Gabaon 21:30 et non praeualuit Dauid ire ad altare ut ibi

for he was afraid because of the sword of the angel of the LORD.

obsecraret Deum nimio enim fuerat timore perterritus uidens gladium angeli Domini

22:1 Then David said, This is the house of the LORD God, 22:1 dixitque Dauid haec est domus Dei et hoc altare in ho-

and this is the altar of the burnt offering for Israel.

22:2 And David commanded to gather together the locaustum Israhel 22:2 et praecepit ut congregarentur om-

strangers that were in the land of Israel; and he set masons to hew wrought stones to build the house of God.

nes proselyti de terra Israhel et constituit ex eis latomos ad caedendos lapides et poliendos ut aedificaretur domus Dei

22:3 And David prepared iron in abundance for the nails 22:3 ferrum quoque plurimum ad clauos ianuarum et ad com-

for the doors of the gates, and for the joinings; and brass in abundance without weight;

missuras atque iuncturas praeparauit Dauid et aeris pondus

22:4 Also cedar trees in abundance: for the Zidonians and innumerabile 22:4 ligna quoque cedrina non poterant aes-

they of Tyre brought much cedar wood to David.

22:5 And David said, Solomon my son is young and ten- timari quae Sidonii et Tyrii deportauerant ad Dauid 22:5 et

der, and the house that is to be builded for the LORD must be exceeding magnifical, of fame and of glory throughout all countries: I will therefore now make preparation for it. So David prepared abundantly before his death.

dixit Dauid Salomon filius meus puer paruulus est et delicatus domus autem quam aedificari uolo Domino talis esse debet ut in cunctis regionibus nominetur praeparabo ergo ei necessaria et ob hanc causam ante mortem suam omnes pa-

22:6 Then he called for Solomon his son, and charged him rauit inpensas 22:6 uocauitque Salomonem filium suum et

to build an house for the LORD God of Israel.

22:7 And David said to Solomon, My son, as for me, it praecepit ei ut aedificaret domum Domino Deo Israhel 22:7

was in my mind to build an house unto the name of the LORD my God:

dixitque Dauid ad Salomonem fili mi uoluntatis meae fuit ut

22:8 But the word of the LORD came to me, saying, Thou aedificarem domum nomini Domini Dei mei 22:8 sed fac-

hast shed blood abundantly, and hast made great wars: thou shalt not build an house unto my name, because thou

hast shed much blood upon the earth in my sight. tus est ad me sermo Domini dicens multum sanguinem effu-

disti et plurima bella bellasti non poteris aedificare domum

22:9 Behold, a son shall be born to thee, who shall be a nomini meo tanto effuso sanguine coram me 22:9 filius qui

man of rest; and I will give him rest from all his enemies round about: for his name shall be Solomon, and I will

give peace and quietness unto Israel in his days. nascetur tibi et erit uir quietissimus faciam enim eum requiescere ab omnibus inimicis suis per circuitum et ob hanc causam pacificus uocabitur et pacem et otium dabo in Israhel

22:10 He shall build an house for my name; and he shall cunctis diebus eius 22:10 ipse aedificabit domum nomini

be my son, and I will be his father; and I will establish the throne of his kingdom over Israel for ever.

meo et ipse erit mihi in filium et ego ero ei in patrem fir-

22:11 Now, my son, the LORD be with thee; and prosper maboque solium regni eius super Israhel in aeternum 22:11

thou, and build the house of the LORD thy God, as he hath said of thee.

nunc ergo fili mi sit Dominus tecum et prosperare et aedifica

22:12 Only the LORD give thee wisdom and understand- domum Domino Deo tuo sicut locutus est de te 22:12 det ing, and give thee charge concerning Israel, that thou mayest keep the law of the LORD thy God.

quoque tibi Dominus prudentiam et sensum ut regere possis

22:13 Then shalt thou prosper, if thou takest heed to ful- Israhel et custodire legem Domini Dei tui 22:13 tunc enim fil the statutes and judgments which the LORD charged Moses with concerning Israel: be strong, and of good

courage; dread not, nor be dismayed. proficere poteris si custodieris mandata et iudicia quae prae-

Liber I Paralipomenon

cepit Dominus Mosi ut doceret Israhel confortare uiriliter age ne timeas neque paueas 22:14 ecce ego in paupertatula 22:14 Now, behold, in my trouble I have prepared for the

house of the LORD an hundred thousand talents of gold, and a thousand thousand talents of silver; and of brass and iron without weight; for it is in abundance: timber also and stone have I prepared; and thou mayest add thereto.

mea praeparaui inpensas domus Domini auri talenta centum milia et argenti mille milia talentorum aeris uero et ferri non est pondus uincitur enim numerus magnitudine ligna et lapides praeparaui ad uniuersa inpendia 22:15 habes quoque 22:15 Moreover there are workmen with thee in abun-

dance, hewers and workers of stone and timber, and all manner of cunning men for every manner of work.

plurimos artifices latomos et cementarios artificesque lignorum et omnium artium ad faciendum opus prudentissimos 22:16 in auro et argento aere et ferro cuius non est numerus 22:16 Of the gold, the silver, and the brass, and the iron,

there is no number. Arise therefore, and be doing, and the LORD be with thee.

surge igitur et fac et erit Dominus tecum 22:17 praecepit

22:17 David also commanded all the princes of Israel to help Solomon his son, saying,

quoque Dauid cunctis principibus Israhel ut adiuuarent Salomonem filium suum 22:18 cernitis inquiens quod Domi- 22:18 Is not the LORD your God with you? and hath he

not given you rest on every side? for he hath given the inhabitants of the land into mine hand; and the land is sub-

nus Deus uester uobiscum sit et dederit uobis requiem per dued before the LORD, and before his people. circuitum et tradiderit omnes inimicos in manu uestra et subiecta sit terra coram Domino et coram populo eius 22:19 22:19 Now set your heart and your soul to seek the LORD

your God; arise therefore, and build ye the sanctuary of the LORD God, to bring the ark of the covenant of the LORD, and the holy vessels of God, into the house that is to be built to the name of the LORD.

praebete igitur corda uestra et animas uestras ut quaeratis Dominum Deum uestrum et consurgite et aedificate sanctuarium Domino Deo ut introducatur arca foederis Domini et uasa Domino consecrata in domum quae aedificatur nomini Domini

23:1 igitur Dauid senex et plenus dierum regem constituit 23:1 So when David was old and full of days, he made

Solomon his son king over Israel.

Salomonem filium suum super Israhel 23:2 et congregauit 23:2 And he gathered together all the princes of Israel,

with the priests and the Levites.

omnes principes Israhel et sacerdotes atque Leuitas 23:3 23:3 Now the Levites were numbered from the age of

thirty years and upward: and their number by their polls, man by man, was thirty and eight thousand.

numeratique sunt Leuitae a triginta annis et supra et inuenta sunt triginta octo milia uirorum 23:4 ex his electi sunt et di- 23:4 Of which, twenty and four thousand were to set for-

ward the work of the house of the LORD; and six thousand were officers and judges:

stributi in ministerium domus Domini uiginti quattuor milia praepositorum autem et iudicum sex milia 23:5 porro quat- 23:5 Moreover four thousand were porters; and four thou-

sand praised the LORD with the instruments which I made, said David, to praise therewith.

tuor milia ianitores et totidem psaltae canentes Domino in organis quae fecerat ad canendum 23:6 et distribuit eos Da- 23:6 And David divided them into courses among the sons

of Levi, namely, Gershon, Kohath, and Merari.

uid per uices filiorum Leui Gersom uidelicet et Caath et Merari 23:7 Gersom Leedan et Semei 23:8 filii Leedan prin- 23:7 Of the Gershonites were, Laadan, and Shimei.

23:8 The sons of Laadan; the chief was Jehiel, and Zetham, and Joel, three.

ceps Ieihel et Zetham et Iohel tres 23:9 filii Semei Salo-

23:9 The sons of Shimei; Shelomith, and Haziel, and Haran, three. These were the chief of the fathers of Laadan.

mith et Ozihel et Aran tres isti principes familiarum Leedan

Biblia Sacra Vulgata

23:10 And the sons of Shimei were, Jahath, Zina, and 23:10 porro filii Semei Ieeth et Ziza et Iaus et Baria isti filii

Jeush, and Beriah. These four were the sons of Shimei.

23:11 And Jahath was the chief, and Zizah the second: but Semei quattuor 23:11 erat autem Ieeth prior Ziza secundus

Jeush and Beriah had not many sons; therefore they were in one reckoning, according to their father’s house.

porro Iaus et Baria non habuerunt plurimos filios et idcirco

23:12 The sons of Kohath; Amram, Izhar, Hebron, and in una familia unaque domo conputati sunt 23:12 filii Caath

Uzziel, four.

23:13 The sons of Amram; Aaron and Moses: and Aaron Amram et Isaar Hebron et Ozihel quattuor 23:13 filii Am-

was separated, that he should sanctify the most holy things, he and his sons for ever, to burn incense before the LORD, to minister unto him, and to bless in his name for ever.

ram Aaron et Moses separatusque est Aaron ut ministraret in sancto sanctorum ipse et filii eius in sempiternum et adoleret incensum Domino secundum ritum suum ac benedice-

23:14 Now concerning Moses the man of God, his sons ret nomini eius in perpetuum 23:14 Mosi quoque hominis

were named of the tribe of Levi.

23:15 The sons of Moses were, Gershom, and Eliezer. Dei filii adnumerati sunt in tribu Leui 23:15 filii Mosi Ger-23:16 Of the sons of Gershom, Shebuel was the chief. som et Eliezer 23:16 filii Gersom Subuhel primus 23:17

23:17 And the sons of Eliezer were, Rehabiah the chief. And Eliezer had none other sons; but the sons of Rehabiah were very many.

fuerunt autem filii Eliezer Roobia primus et non erant Elie-

23:18 Of the sons of Izhar; Shelomith the chief. zer filii alii porro filii Roobia multiplicati sunt nimis 23:18 23:19 Of the sons of Hebron; Jeriah the first, Amariah the filii Isaar Salumith primus 23:19 filii Hebron Ieriau primus

second, Jahaziel the third, and Jekameam the fourth.

23:20 Of the sons of Uzziel; Micah the first and Jesiah the Amarias secundus Iazihel tertius Iecmaam quartus 23:20 fi-

second.

23:21 The sons of Merari; Mahli, and Mushi. The sons of lii Ozihel Micha primus Iesia secundus 23:21 filii Merari

Mahli; Eleazar, and Kish.

23:22 And Eleazar died, and had no sons, but daughters: Mooli et Musi filii Mooli Eleazar et Cis 23:22 mortuus est

and their brethren the sons of Kish took them.

autem Eleazar et non habuit filios sed filias acceperuntque

23:23 The sons of Mushi; Mahli, and Eder, and Jeremoth, eas filii Cis fratres earum 23:23 filii Musi Mooli et Eder et

three.

23:24 These were the sons of Levi after the house of their Ierimuth tres 23:24 hii filii Leui in cognationibus et fami-

fathers; even the chief of the fathers, as they were counted by number of names by their polls, that did the work for the service of the house of the LORD, from the age of twenty years and upward.

liis suis principes per uices et numerum capitum singulorum qui faciebant opera ministerii domus Domini a uiginti

23:25 For David said, The LORD God of Israel hath given annis et supra 23:25 dixit enim Dauid requiem dedit Do-

rest unto his people, that they may dwell in Jerusalem for ever:

minus Deus Israhel populo suo et habitationem Hierusalem

23:26 And also unto the Levites; they shall no more carry usque in aeternum 23:26 nec erit officii Leuitarum ut ultra

the tabernacle, nor any vessels of it for the service thereof.

portent tabernaculum et omnia uasa eius ad ministrandum

23:27 For by the last words of David the Levites were 23:27 iuxta praecepta quoque Dauid nouissima supputabitur

numbered from twenty years old and above:

23:28 Because their office was to wait on the sons of numerus filiorum Leui a uiginti annis et supra 23:28 et erunt

Aaron for the service of the house of the LORD, in the courts, and in the chambers, and in the purifying of all holy things, and the work of the service of the house of God;

sub manu filiorum Aaron in cultum domus Domini in uestibulis et in exedris et in loco purificationis et in sanctuario

23:29 Both for the shewbread, and for the fine flour for et in uniuersis operibus ministerii templi Domini 23:29 sa-

meat offering, and for the unleavened cakes, and for that which is baked in the pan, and for that which is fried, and

for all manner of measure and size; cerdotes autem super panes propositionis et ad similae sacrificium et ad lagana et azyma et sartaginem et ad feruentem

Liber I Paralipomenon

similam et super omne pondus atque mensuram 23:30 Leui- 23:30 And to stand every morning to thank and praise the

LORD, and likewise at even:

tae uero ut stent mane ad confitendum et canendum Domino similiterque ad uesperam 23:31 tam in oblatione holocau- 23:31 And to offer all burnt sacrifices unto the LORD in

the sabbaths, in the new moons, and on the set feasts, by number, according to the order commanded unto them,

storum Domini quam in sabbatis et kalendis et sollemnita- continually before the LORD: tibus reliquis iuxta numerum et caerimonias uniuscuiusque rei iugiter coram Domino 23:32 et custodiant obseruationes 23:32 And that they should keep the charge of the taberna-

cle of the congregation, and the charge of the holy place, and the charge of the sons of Aaron their brethren, in the

tabernaculi foederis et ritum sanctuarii et obseruationem fi- service of the house of the LORD. liorum Aaron fratrum suorum ut ministrent in domo Domini

24:1 porro filiis Aaron hae partitiones erunt filii Aaron Na- 24:1 Now these are the divisions of the sons of Aaron. The

sons of Aaron; Nadab, and Abihu, Eleazar, and Ithamar.

dab et Abiu et Eleazar et Ithamar 24:2 mortui sunt autem 24:2 But Nadab and Abihu died before their father, and

had no children: therefore Eleazar and Ithamar executed the priest’s office.

Nadab et Abiu ante patrem suum absque liberis sacerdotioque functus est Eleazar et Ithamar 24:3 et diuisit eos Da- 24:3 And David distributed them, both Zadok of the sons

of Eleazar, and Ahimelech of the sons of Ithamar, according to their offices in their service.

uid id est Sadoc de filiis Eleazar et Ahimelech de filiis Ithamar secundum uices suas et ministerium 24:4 inuentique 24:4 And there were more chief men found of the sons of

Eleazar than of the sons of Ithamar, and thus were they divided. Among the sons of Eleazar there were sixteen chief men of the house of their fathers, and eight among the sons of Ithamar according to the house of their fathers.

sunt multo plures filii Eleazar in principibus uiris quam filii Ithamar diuisit autem eis hoc est filiis Eleazar principes per familias sedecim et filiis Ithamar per familias et domos suas octo 24:5 porro diuisit utrasque inter se familias sor- 24:5 Thus were they divided by lot, one sort with another;

for the governors of the sanctuary, and governors of the house of God, were of the sons of Eleazar, and of the sons

tibus erant enim principes sanctuarii et principes Dei tam of Ithamar. de filiis Eleazar quam de filiis Ithamar 24:6 descripsitque 24:6 And Shemaiah the son of Nethaneel the scribe, one of

the Levites, wrote them before the king, and the princes, and Zadok the priest, and Ahimelech the son of Abiathar, and before the chief of the fathers of the priests and Levites: one principal household being taken for Eleazar, and one taken for Ithamar.

eos Semeias filius Nathanahel scriba Leuites coram rege et principibus et Sadoc sacerdote et Ahimelech filio Abiathar principibus quoque familiarum sacerdotalium et leuiticarum unam domum quae ceteris praeerat Eleazar et al-

teram domum quae sub se habebat ceteros Ithamar 24:7 24:7 Now the first lot came forth to Jehoiarib, the second

to Jedaiah,

exiuit autem sors prima Ioiarib secunda Iedeiae 24:8 ter- 24:8 The third to Harim, the fourth to Seorim, tia Arim quarta Seorim 24:9 quinta Melchia sexta Maiman 24:9 The fifth to Malchijah, the sixth to Mijamin, 24:10 septima Accos octaua Abia 24:11 nona Hiesu decima 24:10 The seventh to Hakkoz, the eighth to Abijah,

24:11 The ninth to Jeshuah, the tenth to Shecaniah,

Sechenia 24:12 undecima Eliasib duodecima Iacim 24:13 _{24:12 The eleventh to Eliashib, the twelfth to Jakim,}

24:13 The thirteenth to Huppah, the fourteenth to Jeshebeab,

tertiadecima Oppa quartadecima Isbaal 24:14 quintadecima

24:14 The fifteenth to Bilgah, the sixteenth to Immer,

Belga sextadecima Emmer 24:15 septimadecima Ezir octa- 24:15 The seventeenth to Hezir, the eighteenth to Aphses,

uadecima Hapses 24:16 nonadecima Phetheia uicesima Ie- 24:16 The nineteenth to Pethahiah, the twentieth to Je-

hezekel,

zecel 24:17 uicesima prima Iachin uicesima secunda Ga- 24:17 The one and twentieth to Jachin, the two and twen-

tieth to Gamul,

Biblia Sacra Vulgata

24:18 The three and twentieth to Delaiah, the four and mul 24:18 uicesima tertia Dalaiau uicesima quarta Mazziau

twentieth to Maaziah.

24:19 These were the orderings of them in their service 24:19 hae uices eorum secundum ministeria sua ut ingredian-

to come into the house of the LORD, according to their manner, under Aaron their father, as the LORD God of

Israel had commanded him. tur domum Domini et iuxta ritum suum sub manu Aaron pat-24:20 And the rest of the sons of Levi were these: Of ris eorum sicut praecepit Dominus Deus Israhel 24:20 porro

the sons of Amram; Shubael: of the sons of Shubael; Jehdeiah.

filiorum Leui qui reliqui fuerant de filiis Amram erat Suba-

24:21 Concerning Rehabiah: of the sons of Rehabiah, the hel et filiis Subahel Iedeia 24:21 de filiis quoque Roobiae

first was Isshiah.

24:22 Of the Izharites; Shelomoth: of the sons of Shelo- princeps Iesias 24:22 Isaaris uero Salemoth filiusque Sale-

moth; Jahath.

24:23 And the sons of Hebron; Jeriah the first, Amariah moth Iaath 24:23 filiusque eius Ieriahu Amarias secundus

the second, Jahaziel the third, Jekameam the fourth.

24:24 Of the sons of Uzziel; Michah: of the sons of Iazihel tertius Iecmaam quartus 24:24 filius Ozihel Micha

Michah; Shamir.

24:25 The brother of Michah was Isshiah: of the sons of filius Micha Samir 24:25 frater Micha Iesia filiusque Ie-

Isshiah; Zechariah.

24:26 The sons of Merari were Mahli and Mushi: the sons siae Zaccharias 24:26 filii Merari Mooli et Musi filius Io-

of Jaaziah; Beno. 24:27 The sons of Merari by Jaaziah; Beno, and Shoham, ziau Benno 24:27 filius quoque Merari Oziau et Soem et and Zaccur, and Ibri.

24:28 Of Mahli came Eleazar, who had no sons. Zacchur et Hebri 24:28 porro Mooli filius Eleazar qui non 24:29 Concerning Kish: the son of Kish was Jerahmeel. habebat liberos 24:29 filius uero Cis Ierahemel 24:30 filii

24:30 The sons also of Mushi; Mahli, and Eder, and Jerimoth. These were the sons of the Levites after the house of their fathers.

Musi Mooli Eder et Ierimoth isti filii Leui secundum domos

24:31 These likewise cast lots over against their brethren familiarum suarum 24:31 miseruntque et ipsi sortes contra the sons of Aaron in the presence of David the king, and Zadok, and Ahimelech, and the chief of the fathers of the priests and Levites, even the principal fathers over against their younger brethren.

fratres suos filios Aaron coram Dauid rege et Sadoc et Ahimelech et principibus familiarum sacerdotalium et leuiticarum tam maiores quam minores omnes sors aequaliter diuidebat

25:1 Moreover David and the captains of the host sepa- 25:1 igitur Dauid et magistratus exercitus secreuerunt in

rated to the service of the sons of Asaph, and of Heman, and of Jeduthun, who should prophesy with harps, with psalteries, and with cymbals: and the number of the workmen according to their service was:

ministerium filios Asaph et Heman et Idithun qui prophetarent in citharis et psalteriis et cymbalis secundum nume-

25:2 Of the sons of Asaph; Zaccur, and Joseph, and Netha- rum suum dedicato sibi officio seruientes 25:2 de filiis As-

niah, and Asarelah, the sons of Asaph under the hands of Asaph, which prophesied according to the order of the

king. aph Zacchur et Ioseph et Nathania et Asarela filii Asaph sub 25:3 Of Jeduthun: the sons of Jeduthun; Gedaliah, and manu Asaph prophetantis iuxta regem 25:3 porro Idithun

Zeri, and Jeshaiah, Hashabiah, and Mattithiah, six, under the hands of their father Jeduthun, who prophesied with a

harp, to give thanks and to praise the LORD. filii Idithun Godolias Sori Iesaias et Sabias et Matthathias

sex sub manu patris sui Idithun qui in cithara prophetabat su-

25:4 Of Heman: the sons of Heman: Bukkiah, Matta- per confitentes et laudantes Dominum 25:4 Heman quoque

niah, Uzziel, Shebuel, and Jerimoth, Hananiah, Hanani, Eliathah, Giddalti, and Romamtiezer, Joshbekashah, Mal-

lothi, Hothir, and Mahazioth: filii Heman Bocciau Matthaniau Ozihel Subuhel et Ierimoth

Ananias Anani Elietha Geddelthi et Romemthiezer et Ies-

25:5 All these were the sons of Heman the king’s seer in bacasa Mellothi Othir Mazioth 25:5 omnes isti filii Heman

the words of God, to lift up the horn. And God gave to Heman fourteen sons and three daughters.

uidentis regis in sermonibus Dei ut exaltaret cornu deditque Deus Heman filios quattuordecim et filias tres 25:6 uniuersi 25:6 All these were under the hands of their father for song

in the house of the LORD, with cymbals, psalteries, and harps, for the service of the house of God, according to the

sub manu patris sui ad cantandum in templo Domini distri- king’s order to Asaph, Jeduthun, and Heman. buti erant in cymbalis et psalteriis et citharis in ministeria domus Domini iuxta regem Asaph uidelicet et Idithun et Heman 25:7 fuit autem numerus eorum cum fratribus suis qui 25:7 So the number of them, with their brethren that were

instructed in the songs of the LORD, even all that were cunning, was two hundred fourscore and eight.

erudiebant canticum Domini cuncti doctores ducenti octoginta octo 25:8 miseruntque sortes per uices suas ex aequo 25:8 And they cast lots, ward against ward, as well the

small as the great, the teacher as the scholar.

tam maior quam minor doctus pariter et indoctus 25:9 eg- 25:9 Now the first lot came forth for Asaph to Joseph: the

second to Gedaliah, who with his brethren and sons were twelve:

ressaque est sors prima Ioseph qui erat de Asaph secunda Godoliae ipsi et filiis eius et fratribus duodecim 25:10 ter- 25:10 The third to Zaccur, he, his sons, and his brethren,

were twelve:

tia Zacchur filiis et fratribus eius duodecim 25:11 quarta Is- 25:11 The fourth to Izri, he, his sons, and his brethren,

were twelve:

ari filiis et fratribus eius duodecim 25:12 quinta Nathaniae 25:12 The fifth to Nethaniah, he, his sons, and his

brethren, were twelve:

filiis et fratribus eius duodecim 25:13 sexta Bocciau filiis 25:13 The sixth to Bukkiah, he, his sons, and his brethren,

were twelve:

et fratribus eius duodecim 25:14 septima Israhela filiis et 25:14 The seventh to Jesharelah, he, his sons, and his

brethren, were twelve:

fratribus eius duodecim 25:15 octaua Isaiae filiis et fratri- 25:15 The eighth to Jeshaiah, he, his sons, and his

brethren, were twelve:

bus eius duodecim 25:16 nona Matthaniae filiis et fratribus 25:16 The ninth to Mattaniah, he, his sons, and his

brethren, were twelve:

eius duodecim 25:17 decima Semeiae filiis et fratribus eius 25:17 The tenth to Shimei, he, his sons, and his brethren,

were twelve:

duodecim 25:18 undecima Ezrahel filiis et fratribus eius 25:18 The eleventh to Azareel, he, his sons, and his

brethren, were twelve:

duodecim 25:19 duodecima Asabiae filiis et fratribus eius 25:19 The twelfth to Hashabiah, he, his sons, and his

brethren, were twelve:

duodecim 25:20 tertiadecima Subahel filiis et fratribus eius 25:20 The thirteenth to Shubael, he, his sons, and his

brethren, were twelve:

duodecim 25:21 quartadecima Matthathiae filiis et fratribus 25:21 The fourteenth to Mattithiah, he, his sons, and his

brethren, were twelve:

eius duodecim 25:22 quintadecima Ierimoth filiis et fratri- 25:22 The fifteenth to Jeremoth, he, his sons, and his

brethren, were twelve:

bus eius duodecim 25:23 sextadecima Ananiae filiis et frat- 25:23 The sixteenth to Hananiah, he, his sons, and his

brethren, were twelve:

ribus eius duodecim 25:24 septimadecima Iesbocasae filiis 25:24 The seventeenth to Joshbekashah, he, his sons, and

his brethren, were twelve:

et fratribus eius duodecim 25:25 octauadecima Anani filiis 25:25 The eighteenth to Hanani, he, his sons, and his

brethren, were twelve:

et fratribus eius duodecim 25:26 nonadecima Mellothi filiis 25:26 The nineteenth to Mallothi, he, his sons, and his

brethren, were twelve:

et fratribus eius duodecim 25:27 uicesima Eliatha filiis et 25:27 The twentieth to Eliathah, he, his sons, and his

brethren, were twelve:

fratribus eius duodecim 25:28 uicesima prima Othir filiis et 25:28 The one and twentieth to Hothir, he, his sons, and

his brethren, were twelve:

fratribus eius duodecim 25:29 uicesima secunda Godollathi 25:29 The two and twentieth to Giddalti, he, his sons, and

his brethren, were twelve:

filiis et fratribus eius duodecim 25:30 uicesima tertia Ma- 25:30 The three and twentieth to Mahazioth, he, his sons,

and his brethren, were twelve:

ziuth filiis et fratribus eius duodecim 25:31 uicesima quarta 25:31 The four and twentieth to Romamtiezer, he, his

sons, and his brethren, were twelve.

Romamthiezer filiis et fratribus eius duodecim

26:1 diuisiones autem ianitorum de Coritis Mesellemia fi- 26:1 Concerning the divisions of the porters: Of the Ko-

rhites was Meshelemiah the son of Kore, of the sons of Asaph.

Biblia Sacra Vulgata

26:2 And the sons of Meshelemiah were, Zechariah the lius Core de filiis Asaph 26:2 filii Mesellemiae Zaccharias

firstborn, Jediael the second, Zebadiah the third, Jathniel the fourth,

primogenitus Iadihel secundus Zabadias tertius Iathanahel

26:3 Elam the fifth, Jehohanan the sixth, Elioenai the sev- quartus 26:3 Ahilam quintus Iohanan sextus Helioenai seenth.

26:4 Moreover the sons of Obededom were, Shemaiah the ptimus 26:4 filii autem Obededom Semeias primogenitus

firstborn, Jehozabad the second, Joah the third, and Sacar the fourth, and Nethaneel the fifth.

Iozabad secundus Iohaa tertius Sachar quartus Nathanahel

26:5 Ammiel the sixth, Issachar the seventh, Peulthai the quintus 26:5 Amihel sextus Isachar septimus Phollathi oc-

eighth: for God blessed him.

26:6 Also unto Shemaiah his son were sons born, that tauus quia benedixit illi Dominus 26:6 Semeiae autem filio

ruled throughout the house of their father: for they were mighty men of valour.

eius nati sunt filii praefecti familiarum suarum erant enim

26:7 The sons of Shemaiah; Othni, and Rephael, and uiri fortissimi 26:7 filii ergo Semeiae Othni et Raphahel et

Obed, Elzabad, whose brethren were strong men, Elihu, and Semachiah.

Obedihel Zabad fratres eius uiri fortissimi Heliu quoque et

26:8 All these of the sons of Obededom: they and their Samachias 26:8 omnes hii de filiis Obededom ipsi et filii et

sons and their brethren, able men for strength for the service, were threescore and two of Obededom.

fratres eorum fortissimi ad ministrandum sexaginta duo de

26:9 And Meshelemiah had sons and brethren, strong Obededom 26:9 porro Mesellamiae filii et fratres robustis-

men, eighteen.

26:10 Also Hosah, of the children of Merari, had sons; simi decem et octo 26:10 de Hosa autem id est de filiis Me-

Simri the chief, (for though he was not the firstborn, yet his father made him the chief;)

rari Semri princeps non enim habuerat primogenitum et id-

26:11 Hilkiah the second, Tebaliah the third, Zechariah the circo posuerat eum pater eius in principem 26:11 Helchias fourth: all the sons and brethren of Hosah were thirteen.

secundus Tabelias tertius Zaccharias quartus omnes hii filii

26:12 Among these were the divisions of the porters, even et fratres Hosa tredecim 26:12 hii diuisi sunt in ianitores ut

among the chief men, having wards one against another, to minister in the house of the LORD.

semper principes custodiarum sicut et fratres eorum mini-

26:13 And they cast lots, as well the small as the great, strarent in domo Domini 26:13 missae sunt autem sortes ex

according to the house of their fathers, for every gate.

aequo et paruis et magnis per familias suas in unamquam-

26:14 And the lot eastward fell to Shelemiah. Then for que portarum 26:14 cecidit igitur sors orientalis Selemiae

Zechariah his son, a wise counsellor, they cast lots; and his lot came out northward.

porro Zacchariae filio eius uiro prudentissimo et erudito sor-

26:15 To Obededom southward; and to his sons the house tito obtigit plaga septentrionalis 26:15 Obededom uero et

of Asuppim.

filiis eius ad austrum in qua parte domus erat seniorum con-

26:16 To Shuppim and Hosah the lot came forth westward, cilium 26:16 Sepphima et Hosa ad occidentem iuxta portam

with the gate Shallecheth, by the causeway of the going up, ward against ward.

quae ducit ad uiam ascensionis custodia contra custodiam

26:17 Eastward were six Levites, northward four a day, 26:17 ad orientem uero Leuitae sex et ad aquilonem quat-

southward four a day, and toward Asuppim two and two.

tuor per diem atque ad meridiem similiter in die quattuor et

26:18 At Parbar westward, four at the causeway, and two ubi erat concilium bini et bini 26:18 in cellulis quoque ia-

at Parbar.

nitorum ad occidentem quattuor in uia binique per cellulas

26:19 These are the divisions of the porters among the 26:19 hae sunt diuisiones ianitorum filiorum Core et Merari

sons of Kore, and among the sons of Merari. 26:20 And of the Levites, Ahijah was over the treasures of 26:20 porro Achias erat super thesauros domus Dei ac uasa the house of God, and over the treasures of the dedicated things.

Liber I Paralipomenon

sanctorum 26:21 filii Ledan filii Gersonni de Ledan prin- 26:21 As concerning the sons of Laadan; the sons of the

Gershonite Laadan, chief fathers, even of Laadan the Gershonite, were Jehieli.

cipes familiarum Ledan et Gersonni Ieiheli 26:22 filii Iei-

26:22 The sons of Jehieli; Zetham, and Joel his brother, which were over the treasures of the house of the LORD.

heli Zathan et Iohel frater eius super thesauros domus Domini 26:23 Amramitis et Isaaritis et Hebronitis et Oziheli- 26:23 Of the Amramites, and the Izharites, the Hebronites,

and the Uzzielites:

tibus 26:24 Subahel autem filius Gersom filii Mosi praepo- 26:24 And Shebuel the son of Gershom, the son of Moses,

was ruler of the treasures.

situs thesauris 26:25 fratres quoque eius Eliezer cuius filius 26:25 And his brethren by Eliezer; Rehabiah his son, and

Jeshaiah his son, and Joram his son, and Zichri his son, and Shelomith his son.

Raabia et huius filius Isaias et huius filius Ioram huius quoque filius Zechri sed et huius filius Selemith 26:26 ipse Se- 26:26 Which Shelomith and his brethren were over all the

treasures of the dedicated things, which David the king, and the chief fathers, the captains over thousands and hun-

lemith et fratres eius super thesauros sanctorum quae sancti- dreds, and the captains of the host, had dedicated. ficauit Dauid rex et principes familiarum et tribuni et centuriones et duces exercitus 26:27 de bellis et manubiis proe- 26:27 Out of the spoils won in battles did they dedicate to

maintain the house of the LORD.

liorum quae consecrauerant ad instaurationem et supellectilem templi Domini 26:28 haec autem uniuersa sanctificauit 26:28 And all that Samuel the seer, and Saul the son of

Kish, and Abner the son of Ner, and Joab the son of Zeruiah, had dedicated; and whosoever had dedicated any thing, it was under the hand of Shelomith, and of his brethren.

Samuhel uidens et Saul filius Cis et Abner filius Ner et Ioab filius Saruiae omnes qui sanctificauerunt ea per manum Salemith et fratrum eius 26:29 Saaritis uero praeerat Chone- 26:29 Of the Izharites, Chenaniah and his sons were for

the outward business over Israel, for officers and judges.

nias et filii eius ad opera forinsecus super Israhel ad docen-

dum et ad iudicandum eos 26:30 porro de Hebronitis Asa- 26:30 And of the Hebronites, Hashabiah and his brethren, men of valour, a thousand and seven hundred, were officers among them of Israel on this side Jordan westward in all the business of the LORD, and in the service of the king.

bias et fratres eius uiri fortissimi mille septingenti praeerant Israheli trans Iordanem contra occidentem in cunctis operibus Domini et in ministerium regis 26:31 Hebronitarum 26:31 Among the Hebronites was Jerijah the chief, even

among the Hebronites, according to the generations of his fathers. In the fortieth year of the reign of David they were sought for, and there were found among them mighty men of valour at Jazer of Gilead.

autem princeps fuit Hieria secundum familias et cognationes eorum quadragesimo anno regni Dauid recensiti sunt et inuenti uiri fortissimi in Iazer Galaad 26:32 fratresque eius 26:32 And his brethren, men of valour, were two thousand

and seven hundred chief fathers, whom king David made rulers over the Reubenites, the Gadites, and the half tribe of Manasseh, for every matter pertaining to God, and affairs of the king.

robustioris aetatis duo milia septingenti principes familiarum praeposuit autem eos Dauid rex Rubenitis et Gadditis et dimidio tribus Manasse in omne ministerium Dei et regis

27:1 filii autem Israhel secundum numerum suum princi- 27:1 Now the children of Israel after their number, to wit, the chief fathers and captains of thousands and hundreds, and their officers that served the king in any matter of the courses, which came in and went out month by month throughout all the months of the year, of every course were twenty and four thousand.

pes familiarum tribuni et centuriones et praefecti qui ministrabant regi iuxta turmas suas ingredientes et egredientes per singulos menses in anno uiginti quattuor milibus singuli praeerant 27:2 primae turmae in primo mense Isboam prae- 27:2 Over the first course for the first month was

Jashobeam the son of Zabdiel: and in his course were twenty and four thousand.

erat filius Zabdihel et sub eo uiginti quattuor milia 27:3

27:3 Of the children of Perez was the chief of all the captains of the host for the first month.

Biblia Sacra Vulgata

de filiis Phares princeps cunctorum principum in exercitu

27:4 And over the course of the second month was Dodai mense primo 27:4 secundi mensis habebat turmam Dudi an Ahohite, and of his course was Mikloth also the ruler: in his course likewise were twenty and four thousand.

Ahohites et post se alterum nomine Macelloth qui regebat

27:5 The third captain of the host for the third month was partem exercitus uiginti quattuor milium 27:5 dux quoque

Benaiah the son of Jehoiada, a chief priest: and in his course were twenty and four thousand.

turmae tertiae in mense tertio erat Banaias filius Ioiadae sa-

27:6 This is that Benaiah, who was mighty among the cerdos et in diuisione sua uiginti quattuor milia 27:6 ipse

thirty, and above the thirty: and in his course was Ammizabad his son.

est Banaias fortissimus inter triginta et super triginta prae-

27:7 The fourth captain for the fourth month was Asahel erat autem turmae ipsius Amizabad filius eius 27:7 quartus

the brother of Joab, and Zebadiah his son after him: and in his course were twenty and four thousand.

mense quarto Asahel frater Ioab et Zabadias filius eius post

27:8 The fifth captain for the fifth month was Shamhuth eum et in turma eius uiginti quattuor milia 27:8 quintus

the Izrahite: and in his course were twenty and four thousand.

mense quinto princeps Samaoth Iezarites et in turma eius ui-

27:9 The sixth captain for the sixth month was Ira the son ginti quattuor milia 27:9 sextus mense sexto Hira filius Ac-

of Ikkesh the Tekoite: and in his course were twenty and four thousand.

ces Thecuites et in turma eius uiginti quattuor milia 27:10

27:10 The seventh captain for the seventh month was Helez the Pelonite, of the children of Ephraim: and in his course were twenty and four thousand.

septimus mense septimo Helles Phallonites de filiis Eph-

27:11 The eighth captain for the eighth month was Sibbe- raim et in turma eius uiginti quattuor milia 27:11 octauus

cai the Hushathite, of the Zarhites: and in his course were twenty and four thousand.

mense octauo Sobochai Asothites de stirpe Zarai et in turma

27:12 The ninth captain for the ninth month was Abiezer eius uiginti quattuor milia 27:12 nonus mense nono Abie-

the Anetothite, of the Benjamites: and in his course were twenty and four thousand.

zer Anathothites de filiis Iemini et in turma eius uiginti quat-

27:13 The tenth captain for the tenth month was Maharai tuor milia 27:13 decimus mense decimo Marai et ipse Nethe Netophathite, of the Zarhites: and in his course were twenty and four thousand.

tophathites de stirpe Zarai et in turma eius uiginti quattuor

27:14 The eleventh captain for the eleventh month was Be- milia 27:14 undecimus mense undecimo Banaias Pharathonaiah the Pirathonite, of the children of Ephraim: and in his course were twenty and four thousand.

nites de filiis Ephraim et in turma eius uiginti quattuor mi-

27:15 The twelfth captain for the twelfth month was Hel- lia 27:15 duodecimus mense duodecimo Holdai Netophadai the Netophathite, of Othniel: and in his course were twenty and four thousand.

thites de stirpe Gothonihel et in turma eius uiginti quattuor

27:16 Furthermore over the tribes of Israel: the ruler of milia 27:16 porro tribubus praeerant Israhel Rubenitis dux

the Reubenites was Eliezer the son of Zichri: of the Simeonites, Shephatiah the son of Maachah:

Eliezer filius Zechri Symeonitis dux Saphatias filius Macha

27:17 Of the Levites, Hashabiah the son of Kemuel: of the 27:17 Leuitis Asabias filius Camuhel Aaronitis Sadoc 27:18

Aaronites, Zadok: 27:18 Of Judah, Elihu, one of the brethren of David: of Issachar, Omri the son of Michael:

Iuda Heliu frater Dauid Isachar Amri filius Michahel 27:19

27:19 Of Zebulun, Ishmaiah the son of Obadiah: of Naph-

tali, Jerimoth the son of Azriel: Zabulonitis Iesmaias filius Abdiae Nepthalitibus Ierimoth fi-27:20 Of the children of Ephraim, Hoshea the son of lius Ozihel 27:20 filiis Ephraim Osee filius Ozaziu dimi-

Azaziah: of the half tribe of Manasseh, Joel the son of Pedaiah:

dio tribus Manasse Iohel filius Phadiae 27:21 et dimidio tri-

27:21 Of the half tribe of Manasseh in Gilead, Iddo the son of Zechariah: of Benjamin, Jaasiel the son of Abner:

bus Manasse in Galaad Iaddo filius Zacchariae Beniamin au-

27:22 Of Dan, Azareel the son of Jeroham. These were tem Iasihel filius Abner 27:22 Dan uero Ezrihel filius Hie-

the princes of the tribes of Israel.

Liber I Paralipomenon

roam hii principes filiorum Israhel 27:23 noluit autem Da- 27:23 But David took not the number of them from twenty

years old and under: because the LORD had said he would increase Israel like to the stars of the heavens.

uid numerare eos a uiginti annis inferius quia dixerat Dominus ut multiplicaret Israhel quasi stellas caeli 27:24 Ioab fi- 27:24 Joab the son of Zeruiah began to number, but he

finished not, because there fell wrath for it against Israel; neither was the number put in the account of the chronicles

lius Saruiae coeperat numerare nec conpleuit quia super hoc of king David. ira inruerat in Israhel et idcirco numerus eorum qui fuerant recensiti non est relatus in fastos regis Dauid 27:25 super 27:25 And over the king’s treasures was Azmaveth the

son of Adiel: and over the storehouses in the fields, in the cities, and in the villages, and in the castles, was Je-

thesauros autem regis fuit Azmoth filius Adihel his autem honathan the son of Uzziah: thesauris qui erant in urbibus et in uicis et in turribus praesidebat Ionathan filius Oziae 27:26 operi autem rustico et 27:26 And over them that did the work of the field for

tillage of the ground was Ezri the son of Chelub:

agricolis qui exercebant terram praeerat Ezri filius Chelub 27:27 uinearumque cultoribus Semeias Ramathites cellis au- 27:27 And over the vineyards was Shimei the Ramathite:

over the increase of the vineyards for the wine cellars was Zabdi the Shiphmite:

tem uinariis Zabdias Aphonites 27:28 nam super oliueta et

27:28 And over the olive trees and the sycomore trees that were in the low plains was Baalhanan the Gederite: and over the cellars of oil was Joash:

ficeta quae erant in campestribus Balanan Gaderites super apothecas autem olei Ioas 27:29 porro armentis quae pas- 27:29 And over the herds that fed in Sharon was Shitrai

the Sharonite: and over the herds that were in the valleys was Shaphat the son of Adlai:

cebantur in Sarona praepositus fuit Setrai Saronites et super boues in uallibus Saphat filius Adli 27:30 super camelos 27:30 Over the camels also was Obil the Ishmaelite: and

over the asses was Jehdeiah the Meronothite:

uero Vbil Ismahelites et super asinos Iadias Meronathites 27:31 super oues quoque Iaziz Agarenus omnes hii princi- 27:31 And over the flocks was Jaziz the Hagerite. All

these were the rulers of the substance which was king David’s.

pes substantiae regis Dauid 27:32 Ionathan autem patruus

27:32 Also Jonathan David’s uncle was a counsellor, a wise man, and a scribe: and Jehiel the son of Hachmoni was with the king’s sons:

Dauid consiliarius uir prudens et litteratus ipse et Iaihel filius Achamoni erant cum filiis regis 27:33 Ahitophel etiam 27:33 And Ahithophel was the king’s counsellor: and

Hushai the Archite was the king’s companion:

consiliarius regis et Husi Arachites amicus regis 27:34 post 27:34 And after Ahithophel was Jehoiada the son of Bena-

iah, and Abiathar: and the general of the king’s army was Joab.

Ahitophel fuit Ioiada filius Banaiae et Abiathar princeps autem exercitus regis erat Ioab

28:1 conuocauit igitur Dauid omnes principes Israhel du- 28:1 And David assembled all the princes of Israel, the

princes of the tribes, and the captains of the companies that ministered to the king by course, and the captains over the thousands, and captains over the hundreds, and the stewards over all the substance and possession of the king, and of his sons, with the officers, and with the mighty men, and with all the valiant men, unto Jerusalem.

ces tribuum et praepositos turmarum qui ministrabant regi tribunos quoque et centuriones et qui praeerant substantiae et possessionibus regis filiosque suos cum eunuchis et potentes et robustissimos quosque in exercitu Hierusalem 28:2 cum- 28:2 Then David the king stood up upon his feet, and said,

Hear me, my brethren, and my people: As for me, I had in mine heart to build an house of rest for the ark of the covenant of the LORD, and for the footstool of our God, and had made ready for the building:

que surrexisset rex et stetisset ait audite me fratres mei et populus meus cogitaui ut aedificarem domum in qua requiesceret arca foederis Domini et scabillum pedum Dei nostri et ad aedificandum omnia praeparaui 28:3 Deus autem dixit mihi 28:3 But God said unto me, Thou shalt not build an house

for my name, because thou hast been a man of war, and hast shed blood.

Biblia Sacra Vulgata

non aedificabis domum nomini meo eo quod sis uir bellator

28:4 Howbeit the LORD God of Israel chose me before all et sanguinem fuderis 28:4 sed elegit Dominus Deus Israhel

the house of my father to be king over Israel for ever: for he hath chosen Judah to be the ruler; and of the house of Judah, the house of my father; and among the sons of my father he liked me to make me king over all Israel:

me de uniuersa domo patris mei ut essem rex super Israhel in sempiternum de Iuda enim elegit principes porro de domo Iuda domum patris mei et de filiis patris mei placuit ei ut me

28:5 And of all my sons, (for the LORD hath given me eligeret regem super cunctum Israhel 28:5 sed et de filiis

many sons,) he hath chosen Solomon my son to sit upon the throne of the kingdom of the LORD over Israel.

meis filios enim multos dedit mihi Dominus elegit Salomonem filium meum ut sederet in throno regni Domini super

28:6 And he said unto me, Solomon thy son, he shall build Israhel 28:6 dixitque mihi Salomon filius tuus aedificabit

my house and my courts: for I have chosen him to be my son, and I will be his father.

domum meam et atria mea ipsum enim elegi mihi in filium

28:7 Moreover I will establish his kingdom for ever, if he et ego ero ei in patrem 28:7 et firmabo regnum eius usque be constant to do my commandments and my judgments, as at this day.

in aeternum si perseuerauerit facere praecepta mea et iudicia

28:8 Now therefore in the sight of all Israel the congrega- sicut et hodie 28:8 nunc igitur coram uniuerso coetu Israhel

tion of the LORD, and in the audience of our God, keep and seek for all the commandments of the LORD your God: that ye may possess this good land, and leave it for an inheritance for your children after you for ever.

audiente Deo nostro custodite et perquirite cuncta mandata Domini Dei nostri ut possideatis terram bonam et relinqua-

28:9 And thou, Solomon my son, know thou the God of tis eam filiis uestris post uos usque in sempiternum 28:9

thy father, and serve him with a perfect heart and with a willing mind: for the LORD searcheth all hearts, and understandeth all the imaginations of the thoughts: if thou seek him, he will be found of thee; but if thou forsake him, he will cast thee off for ever.

tu autem Salomon fili mi scito Deum patris tui et serui ei corde perfecto et animo uoluntario omnia enim corda scrutatur Dominus et uniuersas mentium cogitationes intellegit si quaesieris eum inuenies si autem dereliqueris illum proi-

28:10 Take heed now; for the LORD hath chosen thee to ciet te in aeternum 28:10 nunc ergo quia elegit te Dominus

build an house for the sanctuary: be strong, and do it.

28:11 Then David gave to Solomon his son the pattern of ut aedificares domum sanctuarii confortare et perfice 28:11

the porch, and of the houses thereof, and of the treasuries thereof, and of the upper chambers thereof, and of the in-

ner parlours thereof, and of the place of the mercy seat, dedit autem Dauid Salomoni filio suo descriptionem porti-

cus et templi et cellariorum et cenaculi et cubiculorum in ad-

28:12 And the pattern of all that he had by the spirit, of the ytis et domus propitiationis 28:12 necnon et omnium quae courts of the house of the LORD, and of all the chambers round about, of the treasuries of the house of God, and of

the treasuries of the dedicated things: cogitauerat atriorum et exedrarum per circuitum in thesau-28:13 Also for the courses of the priests and the Levites, ros domus Domini et in thesauros sanctorum 28:13 diuisio-

and for all the work of the service of the house of the LORD, and for all the vessels of service in the house of

the LORD. numque sacerdotalium et leuiticarum in omnia opera domus 28:14 He gave of gold by weight for things of gold, for all Domini et in uniuersa uasa ministerii templi Domini 28:14

instruments of all manner of service; silver also for all instruments of silver by weight, for all instruments of every

kind of service: aurum in pondere per singula uasa ministerii argenti quo-

28:15 Even the weight for the candlesticks of gold, and for que pondus pro uasorum ad opera diuersitate 28:15 sed et their lamps of gold, by weight for every candlestick, and for the lamps thereof: and for the candlesticks of silver by weight, both for the candlestick, and also for the lamps thereof, according to the use of every candlestick.

ad candelabra aurea et ad lucernas eorum aurum pro mensura uniuscuiusque candelabri et lucernarum similiter et in

Liber I Paralipomenon

candelabris argenteis et in lucernis eorum pro diuersitate mensurae pondus argenti tradidit 28:16 aurum quoque dedit 28:16 And by weight he gave gold for the tables of shew-

bread, for every table; and likewise silver for the tables of silver:

in mensas propositionis pro diuersitate mensarum similiter et argentum in alias mensas argenteas 28:17 ad fuscinulas 28:17 Also pure gold for the fleshhooks, and the bowls,

and the cups: and for the golden basons he gave gold by weight for every bason; and likewise silver by weight for

quoque et fialas et turibula ex auro purissimo et leunculos every bason of silver: aureos pro qualitate mensurae pondus distribuit in leunculum et leunculum similiter et in leones argenteos diuersum argenti pondus separauit 28:18 altari autem in quo adoletur 28:18 And for the altar of incense refined gold by weight;

and gold for the pattern of the chariot of the cherubims, that spread out their wings, and covered the ark of the

incensum aurum purissimum dedit ut ex ipso fieret simili- covenant of the LORD. tudo quadrigae cherubin extendentium alas et uelantium arcam foederis Domini 28:19 omnia inquit uenerunt scripta 28:19 All this, said David, the LORD made me understand

in writing by his hand upon me, even all the works of this pattern.

manu Domini ad me ut intellegerem uniuersa opera exemplaris 28:20 dixit quoque Dauid Salomoni filio suo uirili- 28:20 And David said to Solomon his son, Be strong and

of good courage, and do it: fear not, nor be dismayed: for the LORD God, even my God, will be with thee; he will not fail thee, nor forsake thee, until thou hast finished all the work for the service of the house of the LORD.

ter age et confortare et fac ne timeas et ne paueas Dominus enim Deus meus tecum erit et non dimittet te nec derelinquet donec perficias omne opus ministerii domus Domini 28:21 28:21 And, behold, the courses of the priests and the

Levites, even they shall be with thee for all the service of the house of God: and there shall be with thee for all manner of workmanship every willing skilful man, for any manner of service: also the princes and all the people will be wholly at thy commandment.

ecce diuisiones sacerdotum et Leuitarum in omne ministerium domus Domini adsistunt tibi et parati sunt et nouerunt tam principes quam populus facere omnia praecepta tua

29:1 locutusque est Dauid rex ad omnem ecclesiam Sa- 29:1 Furthermore David the king said unto all the congregation, Solomon my son, whom alone God hath chosen, is yet young and tender, and the work is great: for the palace

lomonem filium meum unum elegit Deus adhuc puerum et is not for man, but for the LORD God. tenellum opus autem grande est neque enim homini praeparatur habitatio sed Deo 29:2 ego autem totis uiribus meis 29:2 Now I have prepared with all my might for the house

of my God the gold for things to be made of gold, and the silver for things of silver, and the brass for things of brass, the iron for things of iron, and wood for things of wood; onyx stones, and stones to be set, glistering stones, and of divers colours, and all manner of precious stones, and marble stones in abundance.

praeparaui inpensas domus Dei mei aurum ad uasa aurea et argentum in argentea aes in aenea ferrum in ferrea lignum ad lignea lapides onychinos et quasi stibinos et diuersorum colorum omnem pretiosum lapidem et marmor parium abundantissime 29:3 et super haec quae obtuli in domum Dei 29:3 Moreover, because I have set my affection to the

house of my God, I have of mine own proper good, of gold and silver, which I have given to the house of my God, over and above all that I have prepared for the holy house.

mei de peculio meo aurum et argentum do in templum Dei mei exceptis his quae paraui in aedem sanctam 29:4 tria 29:4 Even three thousand talents of gold, of the gold of

Ophir, and seven thousand talents of refined silver, to overlay the walls of the houses withal:

milia talenta auri de auro Ophir et septem milia talentorum argenti probatissimi ad deaurandos parietes templi 29:5 ut 29:5 The gold for things of gold, and the silver for things

of silver, and for all manner of work to be made by the hands of artificers. And who then is willing to consecrate

ubicumque opus est aurum de auro et ubicumque opus est his service this day unto the LORD?

Biblia Sacra Vulgata

argentum argenti opera fiant per manus artificum et si quis sponte offert impleat manum suam hodie et offerat quod

29:6 Then the chief of the fathers and princes of the tribes uoluerit Domino 29:6 polliciti sunt itaque principes fami-

of Israel and the captains of thousands and of hundreds, with the rulers of the king’s work, offered willingly,

liarum et proceres tribuum Israhel tribuni quoque et centu-

29:7 And gave for the service of the house of God of gold riones et principes possessionum regis 29:7 dederuntque in

five thousand talents and ten thousand drams, and of silver ten thousand talents, and of brass eighteen thousand

talents, and one hundred thousand talents of iron. opera domus Dei auri talenta quinque milia et solidos de-

cem milia argenti talenta decem milia et aeris talenta decem

29:8 And they with whom precious stones were found et octo milia ferri quoque centum milia talentorum 29:8

gave them to the treasure of the house of the LORD, by the hand of Jehiel the Gershonite.

et apud quemcumque inuenti sunt lapides dederunt in the-

29:9 Then the people rejoiced, for that they offered will- saurum domus Domini per manum Ieihel Gersonitis 29:9 ingly, because with perfect heart they offered willingly to the LORD: and David the king also rejoiced with great

joy. laetatusque est populus cum uota sponte promitterent quia

corde toto offerebant ea Domino sed et Dauid rex laetatus est

29:10 Wherefore David blessed the LORD before all the gaudio magno 29:10 et benedixit Domino coram uniuersa

congregation: and David said, Blessed be thou, LORD God of Israel our father, for ever and ever.

multitudine et ait benedictus es Domine Deus Israhel patris

29:11 Thine, O LORD is the greatness, and the power, and nostri ab aeterno in aeternum 29:11 tua est Domine magni-

the glory, and the victory, and the majesty: for all that is in the heaven and in the earth is thine; thine is the kingdom,

O LORD, and thou art exalted as head above all. ficentia et potentia et gloria atque uictoria et tibi laus cuncta

enim quae in caelo sunt et in terra tua sunt tuum Domine re-

29:12 Both riches and honour come of thee, and thou gnum et tu es super omnes principes 29:12 tuae diuitiae et

reignest over all; and in thine hand is power and might; and in thine hand it is to make great, and to give strength

unto all. tua est gloria tu dominaris omnium in manu tua uirtus et po-

29:13 Now therefore, our God, we thank thee, and praise tentia in manu tua magnitudo et imperium omnium 29:13 thy glorious name.

nunc igitur Deus noster confitemur tibi et laudamus nomen

29:14 But who am I, and what is my people, that we should tuum inclitum 29:14 quis ego et quis populus meus ut pos-

be able to offer so willingly after this sort? for all things come of thee, and of thine own have we given thee.

simus haec tibi uniuersa promittere tua sunt omnia et quae de

29:15 For we are strangers before thee, and sojourners, as manu tua accepimus dedimus tibi 29:15 peregrini enim suwere all our fathers: our days on the earth are as a shadow, and there is none abiding.

mus coram te et aduenae sicut omnes patres nostri dies nostri

29:16 O LORD our God, all this store that we have pre- quasi umbra super terram et nulla est mora 29:16 Domine

pared to build thee an house for thine holy name cometh of thine hand, and is all thine own.

Deus noster omnis haec copia quam parauimus ut aedificaretur domus nomini sancto tuo de manu tua est et tua sunt

29:17 I know also, my God, that thou triest the heart, and omnia 29:17 scio Deus meus quod probes corda et simplihast pleasure in uprightness. As for me, in the uprightness of mine heart I have willingly offered all these things: and now have I seen with joy thy people, which are present here, to offer willingly unto thee.

citatem diligas unde et ego in simplicitate cordis mei laetus obtuli uniuersa haec et populum tuum qui hic reppertus est

29:18 O LORD God of Abraham, Isaac, and of Israel, uidi cum ingenti gaudio tibi offerre donaria 29:18 Domine

our fathers, keep this for ever in the imagination of the thoughts of the heart of thy people, and prepare their heart

unto thee: Deus Abraham et Isaac et Israhel patrum nostrorum custodi

Liber I Paralipomenon

in aeternum hanc uoluntatem cordis eorum et semper in uenerationem tui mens ista permaneat 29:19 Salomoni quo- 29:19 And give unto Solomon my son a perfect heart,

to keep thy commandments, thy testimonies, and thy statutes, and to do all these things, and to build the palace,

que filio meo da cor perfectum ut custodiat mandata tua te- for the which I have made provision. stimonia tua caerimonias tuas et faciat uniuersa et aedificet aedem cuius inpensas paraui 29:20 praecepit autem Dauid 29:20 And David said to all the congregation, Now bless

the LORD your God. And all the congregation blessed the LORD God of their fathers, and bowed down their heads,

uniuersae ecclesiae benedicite Domino Deo nostro et bene- and worshipped the LORD, and the king. dixit omnis ecclesia Domino Deo patrum suorum et inclinauerunt se et adorauerunt Deum et deinde regem 29:21 29:21 And they sacrificed sacrifices unto the LORD, and

offered burnt offerings unto the LORD, on the morrow after that day, even a thousand bullocks, a thousand rams, and a thousand lambs, with their drink offerings, and sacrifices in abundance for all Israel:

immolaueruntque uictimas Domino et obtulerunt holocausta die sequenti tauros mille arietes mille agnos mille cum libaminibus suis et uniuerso ritu abundantissime in omnem Israhel 29:22 et comederunt et biberunt coram Domino in 29:22 And did eat and drink before the LORD on that day

with great gladness. And they made Solomon the son of David king the second time, and anointed him unto the

die illo cum grandi laetitia et unxerunt secundo Salomonem LORD to be the chief governor, and Zadok to be priest. filium Dauid unxerunt autem Domino in principem et Sadoc in pontificem 29:23 seditque Salomon super solium Domini 29:23 Then Solomon sat on the throne of the LORD as

king instead of David his father, and prospered; and all Israel obeyed him.

in regem pro Dauid patre suo et cunctis placuit et paruit illi omnis Israhel 29:24 sed et uniuersi principes et potentes et 29:24 And all the princes, and the mighty men, and all the

sons likewise of king David, submitted themselves unto Solomon the king.

cuncti filii regis Dauid dederunt manum et subiecti fuerunt Salomoni regi 29:25 magnificauit ergo Dominus Salomo- 29:25 And the LORD magnified Solomon exceedingly in

the sight of all Israel, and bestowed upon him such royal majesty as had not been on any king before him in Israel.

nem super omnem Israhel et dedit illi gloriam regni qualem nullus habuit ante eum rex Israhel 29:26 igitur Dauid filius 29:26 Thus David the son of Jesse reigned over all Israel. Isai regnauit super uniuersum Israhel 29:27 et dies quibus 29:27 And the time that he reigned over Israel was forty

years; seven years reigned he in Hebron, and thirty and three years reigned he in Jerusalem.

regnauit super Israhel fuerunt quadraginta anni in Hebron regnauit septem annis et in Hierusalem triginta tribus 29:28 29:28 And he died in a good old age, full of days, riches,

and honour: and Solomon his son reigned in his stead.

et mortuus est in senectute bona plenus dierum et diuitiis et gloria regnauitque Salomon filius eius pro eo 29:29 gesta 29:29 Now the acts of David the king, first and last, be-

hold, they are written in the book of Samuel the seer, and in the book of Nathan the prophet, and in the book of Gad

autem Dauid regis priora et nouissima scripta sunt in libro the seer, Samuhel uidentis et in libro Nathan prophetae atque in uolumine Gad uidentis 29:30 uniuersique regni eius et fortitudi- 29:30 With all his reign and his might, and the times that

went over him, and over Israel, and over all the kingdoms of the countries.

nis et temporum quae transierunt sub eo siue in Israhel siue in cunctis regnis terrarum

Biblia Sacra Vulgata

DCCXXX

Liber II Paralipomenon

1:1 confortatus est ergo Salomon filius Dauid in regno suo 1:1 And Solomon the son of David was strengthened in

his kingdom, and the LORD his God was with him, and magnified him exceedingly.

et Dominus erat cum eo et magnificauit eum in excelsum 1:2 praecepitque Salomon uniuerso Israheli tribunis et centu- 1:2 Then Solomon spake unto all Israel, to the captains of

thousands and of hundreds, and to the judges, and to every governor in all Israel, the chief of the fathers.

rionibus et ducibus et iudicibus omnis Israhel et principibus familiarum 1:3 et abiit cum uniuersa multitudine in excel- 1:3 So Solomon, and all the congregation with him, went

to the high place that was at Gibeon; for there was the tabernacle of the congregation of God, which Moses the

sum Gabaon ubi erat tabernaculum foederis Dei quod fecit servant of the LORD had made in the wilderness. Moses famulus Dei in solitudine 1:4 arcam autem Dei ad- 1:4 But the ark of God had David brought up from Kir-

jathjearim to the place which David had prepared for it: for he had pitched a tent for it at Jerusalem.

duxerat Dauid de Cariathiarim in locum quem parauerat ei et

ubi fixerat illi tabernaculum hoc est in Hierusalem 1:5 al- 1:5 Moreover the brasen altar, that Bezaleel the son of Uri, the son of Hur, had made, he put before the tabernacle of the LORD: and Solomon and the congregation sought unto

tare quoque aeneum quod fabricatus fuerat Beselehel filius it. Vri filii Vr ibi erat coram tabernaculo Domini quod et requisiuit Salomon et omnis ecclesia 1:6 ascenditque Salomon 1:6 And Solomon went up thither to the brasen altar before

the LORD, which was at the tabernacle of the congregation, and offered a thousand burnt offerings upon it.

ad altare aeneum coram tabernaculo foederis Domini et obtulit in eo mille hostias 1:7 ecce autem in ipsa nocte appa- 1:7 In that night did God appear unto Solomon, and said

unto him, Ask what I shall give thee.

ruit ei Deus dicens postula quod uis ut dem tibi 1:8 dixit- 1:8 And Solomon said unto God, Thou hast shewed great

mercy unto David my father, and hast made me to reign in his stead.

que Salomon Deo tu fecisti cum Dauid patre meo misericordiam magnam et constituisti me regem pro eo 1:9 nunc 1:9 Now, O LORD God, let thy promise unto David my

father be established: for thou hast made me king over a people like the dust of the earth in multitude.

igitur Domine Deus impleatur sermo tuus quem pollicitus es Dauid patri meo tu enim fecisti me regem super populum tuum multum qui tam innumerabilis est quam puluis terrae 1:10 da mihi sapientiam et intellegentiam ut egrediar coram 1:10 Give me now wisdom and knowledge, that I may go

out and come in before this people: for who can judge this thy people, that is so great?

populo tuo et ingrediar quis enim potest hunc populum tuum digne qui tam grandis est iudicare 1:11 dixit autem Deus 1:11 And God said to Solomon, Because this was in thine

heart, and thou hast not asked riches, wealth, or honour, nor the life of thine enemies, neither yet hast asked long life; but hast asked wisdom and knowledge for thyself, that thou mayest judge my people, over whom I have made thee king:

ad Salomonem quia hoc magis placuit cordi tuo et non postulasti diuitias et substantiam et gloriam neque animas eorum qui te oderunt sed nec dies uitae plurimos petisti autem sapientiam et scientiam ut iudicare possis populum meum super quem constitui te regem 1:12 sapientia et scientia data 1:12 Wisdom and knowledge is granted unto thee; and I

will give thee riches, and wealth, and honour, such as none of the kings have had that have been before thee, neither

sunt tibi diuitias autem et substantiam et gloriam dabo tibi shall there any after thee have the like. ita ut nullus in regibus nec ante te nec post te fuerit similis tui 1:13 uenit ergo Salomon ab excelso Gabaon in Hie- 1:13 Then Solomon came from his journey to the high

place that was at Gibeon to Jerusalem, from before the tabernacle of the congregation, and reigned over Israel.

rusalem coram tabernaculo foederis et regnauit super Israhel 1:14 congregauitque sibi currus et equites et facti sunt ei 1:14 And Solomon gathered chariots and horsemen: and

he had a thousand and four hundred chariots, and twelve thousand horsemen, which he placed in the chariot cities, and with the king at Jerusalem.

Biblia Sacra Vulgata

mille quadringenti currus et duodecim milia equitum et fecit eos esse in urbibus quadrigarum et cum rege in Hierusalem

1:15 And the king made silver and gold at Jerusalem as 1:15 praebuitque rex argentum et aurum in Hierusalem quasi

plenteous as stones, and cedar trees made he as the sycomore trees that are in the vale for abundance.

lapides et cedros quasi sycomoros quae nascuntur in cam-

1:16 And Solomon had horses brought out of Egypt, and pestribus multitudine magna 1:16 adducebantur autem ei et

linen yarn: the king’s merchants received the linen yarn at a price.

equi de Aegypto et de Coa a negotiatoribus regis qui ibant

1:17 And they fetched up, and brought forth out of Egypt et coemebant pretio 1:17 quadrigam equorum sescentis ar-

a chariot for six hundred shekels of silver, and an horse for an hundred and fifty: and so brought they out horses for all the kings of the Hittites, and for the kings of Syria, by their means.

genteis et equum centum quinquaginta similiter de uniuersis regnis Cettheorum et a regibus Syriae emptio celebrabatur

2:1 And Solomon determined to build an house for the 2:1 decreuit autem Salomon aedificare domum nomini name of the LORD, and an house for his kingdom.

2:2 And Solomon told out threescore and ten thousand Domini et palatium sibi 2:2 et numerauit septuaginta mi-

men to bear burdens, and fourscore thousand to hew in the mountain, and three thousand and six hundred to oversee

them. lia uirorum portantium umeris et octoginta milia qui caede-

rent lapides in montibus praepositosque eorum tria milia se-

2:3 And Solomon sent to Huram the king of Tyre, saying, scentos 2:3 misit quoque ad Hiram regem Tyri dicens sicut

As thou didst deal with David my father, and didst send him cedars to build him an house to dwell therein, even so

deal with me. egisti cum Dauid patre meo et misisti ei ligna cedrina ut ae-2:4 Behold, I build an house to the name of the LORD my dificaret sibi domum in qua et habitauit 2:4 sic fac mecum

God, to dedicate it to him, and to burn before him sweet incense, and for the continual shewbread, and for the burnt offerings morning and evening, on the sabbaths, and on the new moons, and on the solemn feasts of the LORD our God. This is an ordinance for ever to Israel.

ut aedificem domum nomini Domini Dei mei et consecrem eam ad adolendum incensum coram illo et fumiganda aromata et ad propositionem panum sempiternam et holocaustomata mane et uespere sabbatis quoque et neomeniis et sollemnitatibus Domini Dei nostri in sempiternum quae man-

2:5 And the house which I build is great: for great is our data sunt Israheli 2:5 domus autem quam aedificare cupio

God above all gods.

magna est magnus est enim Deus noster super omnes de-

2:6 But who is able to build him an house, seeing the os 2:6 quis ergo poterit praeualere ut aedificet ei dignam

heaven and heaven of heavens cannot contain him? who am I then, that I should build him an house, save only to

burn sacrifice before him? domum si caelum et caeli caelorum capere eum non queunt

quantus ego sum ut possim ei aedificare domum sed ad hoc

2:7 Send me now therefore a man cunning to work in gold, tantum ut adoleatur incensum coram illo 2:7 mitte igitur

and in silver, and in brass, and in iron, and in purple, and crimson, and blue, and that can skill to grave with the cunning men that are with me in Judah and in Jerusalem, whom David my father did provide.

mihi uirum eruditum qui nouerit operari in auro et argento aere ferro purpura coccino et hyacintho et qui sciat scalpere celata cum his artificibus quos mecum habeo in Iudaea et

2:8 Send me also cedar trees, fir trees, and algum trees, out in Hierusalem quos praeparauit Dauid pater meus 2:8 sed of Lebanon: for I know that thy servants can skill to cut timber in Lebanon; and, behold, my servants shall be with

thy servants, et ligna cedrina mitte mihi et arceuthina et pinea de Libano scio enim quod serui tui nouerint caedere ligna de Libano et

Liber II Paralipomenon

erunt serui mei cum seruis tuis 2:9 ut parentur mihi ligna 2:9 Even to prepare me timber in abundance: for the house

which I am about to build shall be wonderful great.

plurima domus enim quam cupio aedificare magna est nimis et inclita 2:10 praeterea operariis qui caesuri sunt ligna ser- 2:10 And, behold, I will give to thy servants, the hew-

ers that cut timber, twenty thousand measures of beaten wheat, and twenty thousand measures of barley, and twenty thousand baths of wine, and twenty thousand baths of oil.

uis tuis dabo in cibaria tritici choros uiginti milia et hordei choros totidem olei quoque sata uiginti milia 2:11 dixit au- 2:11 Then Huram the king of Tyre answered in writing,

which he sent to Solomon, Because the LORD hath loved his people, he hath made thee king over them.

tem Hiram rex Tyri per litteras quas miserat Salomoni quia dilexit Dominus populum suum idcirco te regnare fecit super eum 2:12 et addidit dicens benedictus Dominus Deus 2:12 Huram said moreover, Blessed be the LORD God

of Israel, that made heaven and earth, who hath given to David the king a wise son, endued with prudence and understanding, that might build an house for the LORD, and an house for his kingdom.

Israhel qui fecit caelum et terram qui dedit Dauid regi filium sapientem et eruditum et sensatum atque prudentem ut aedificaret domum Domino et palatium sibi 2:13 misi ergo 2:13 And now I have sent a cunning man, endued with

understanding, of Huram my father’s,

tibi uirum prudentem et scientissimum Hiram patrem meum 2:14 filium mulieris de filiabus Dan cuius pater Tyrius fuit qui 2:14 The son of a woman of the daughters of Dan, and his

father was a man of Tyre, skilful to work in gold, and in silver, in brass, in iron, in stone, and in timber, in purple, in blue, and in fine linen, and in crimson; also to grave any manner of graving, and to find out every device which shall be put to him, with thy cunning men, and with the cunning men of my lord David thy father.

nouerit operari in auro et argento et aere et ferro et marmore et lignis in purpura quoque et hyacintho et bysso et coccino et qui sciat celare omnem scalpturam et adinuenire prudenter quodcumque in opere necessarium est cum artificibus tuis et cum artificibus domini mei Dauid patris tui 2:15 triticum 2:15 Now therefore the wheat, and the barley, the oil, and

the wine, which my lord hath spoken of, let him send unto his servants:

ergo et hordeum et oleum et uinum quae pollicitus es domine mi mitte seruis tuis 2:16 nos autem caedemus ligna 2:16 And we will cut wood out of Lebanon, as much as

thou shalt need: and we will bring it to thee in floats by sea to Joppa; and thou shalt carry it up to Jerusalem.

de Libano quot necessaria habueris et adplicabimus ea ratibus per mare in Ioppe tuum erit transferre ea in Hierusalem 2:17 numerauit igitur Salomon omnes uiros proselytos qui er- 2:17 And Solomon numbered all the strangers that were

in the land of Israel, after the numbering wherewith David his father had numbered them; and they were found an hundred and fifty thousand and three thousand and six hundred.

ant in terra Israhel post dinumerationem quam dinumerauit Dauid pater eius et inuenti sunt centum quinquaginta milia et tria milia sescenti 2:18 fecitque ex eis septuaginta milia 2:18 And he set threescore and ten thousand of them to

be bearers of burdens, and fourscore thousand to be hewers in the mountain, and three thousand and six hundred

qui umeris onera portarent et octoginta milia qui lapides in overseers to set the people a work. montibus caederent tria milia autem et sescentos praepositos operum populi

3:1 et coepit Salomon aedificare domum Domini in Hie- 3:1 Then Solomon began to build the house of the LORD

at Jerusalem in mount Moriah, where the Lord appeared unto David his father, in the place that David had prepared

rusalem in monte Moria qui demonstratus fuerat Dauid pa- in the threshingfloor of Ornan the Jebusite. tri eius in loco quem parauerat Dauid in area Ornan Iebusei 3:2 coepit autem aedificare mense secundo anno quarto regni 3:2 And he began to build in the second day of the second

month, in the fourth year of his reign.

Biblia Sacra Vulgata

3:3 Now these are the things wherein Solomon was in- sui 3:3 et haec sunt fundamenta quae iecit Salomon ut ae-

structed for the building of the house of God. The length by cubits after the first measure was threescore cubits, and

the breadth twenty cubits. dificaret domum Dei longitudinis cubitos in mensura prima 3:4 And the porch that was in the front of the house, the sexaginta latitudinis cubitos uiginti 3:4 porticum uero ante

length of it was according to the breadth of the house, twenty cubits, and the height was an hundred and twenty:

and he overlaid it within with pure gold. frontem quae tendebatur in longum iuxta mensuram latitudinis domus cubitorum uiginti porro altitudo centum uiginti cubitorum erat et deaurauit eam intrinsecus auro mun-

3:5 And the greater house he cieled with fir tree, which dissimo 3:5 domum quoque maiorem texit tabulis ligneis

he overlaid with fine gold, and set thereon palm trees and chains.

abiegnis et lamminas auri obrizi adfixit per totum scalpsitque in ea palmas et quasi catenulas se inuicem conplectentes

3:6 And he garnished the house with precious stones for 3:6 strauit quoque pauimentum templi pretiosissimo mar-

beauty: and the gold was gold of Parvaim. 3:7 He overlaid also the house, the beams, the posts, and more decore multo 3:7 porro aurum erat probatissimum de the walls thereof, and the doors thereof, with gold; and graved cherubims on the walls.

cuius lamminis texit domum et trabes eius et postes et parie-

3:8 And he made the most holy house, the length whereof tes et ostia et celauit cherubin in parietibus 3:8 fecit quoque

was according to the breadth of the house, twenty cubits, and the breadth thereof twenty cubits: and he overlaid it

with fine gold, amounting to six hundred talents. domum sancti sanctorum longitudinem iuxta latitudinem domus cubitorum uiginti et latitudinem similiter uiginti cubitorum et lamminis aureis texit eam quasi talentis sescentis

3:9 And the weight of the nails was fifty shekels of gold. 3:9 sed et clauos fecit aureos ita ut singuli claui siclos quin-

And he overlaid the upper chambers with gold.

3:10 And in the most holy house he made two cherubims quagenos adpenderent cenacula quoque texit auro 3:10 fe-

of image work, and overlaid them with gold.

cit etiam in domo sancti sanctorum cherubin duo opere sta-

3:11 And the wings of the cherubims were twenty cubits tuario et texit eos auro 3:11 alae cherubin uiginti cubitis ex-

long: one wing of the one cherub was five cubits, reaching to the wall of the house: and the other wing was likewise

five cubits, reaching to the wing of the other cherub. tendebantur ita ut una ala haberet cubitos quinque et tangeret

parietem domus et altera quinque cubitos habens alam tan-

3:12 And one wing of the other cherub was five cubits, geret alterius cherub 3:12 similiter cherub alterius ala quin-

reaching to the wall of the house: and the other wing was five cubits also, joining to the wing of the other cherub.

que habebat cubitos et tangebat parietem et ala eius altera

3:13 The wings of these cherubims spread themselves quinque cubitorum alam cherub alterius contingebat 3:13

forth twenty cubits: and they stood on their feet, and their faces were inward.

igitur alae utriusque cherubin expansae erant et extendebantur per cubitos uiginti ipsi autem stabant erectis pedibus et

3:14 And he made the vail of blue, and purple, and crim- facies eorum uersae erant ad exteriorem domum 3:14 fe-

son, and fine linen, and wrought cherubims thereon.

cit quoque uelum ex hyacintho purpura coccino et bysso et

3:15 Also he made before the house two pillars of thirty intexuit ei cherubin 3:15 ante fores etiam templi duas co-

and five cubits high, and the chapiter that was on the top of each of them was five cubits.

lumnas quae triginta et quinque cubitos habebant altitudi-

3:16 And he made chains, as in the oracle, and put nis porro capita earum quinque cubitorum 3:16 necnon et

them on the heads of the pillars; and made an hundred pomegranates, and put them on the chains.

quasi catenulas in oraculo et superposuit eas capitibus co-

Liber II Paralipomenon

lumnarum malagranata etiam centum quae catenulis interposuit 3:17 ipsas quoque columnas posuit in uestibulo templi 3:17 And he reared up the pillars before the temple, one

on the right hand, and the other on the left; and called the name of that on the right hand Jachin, and the name of that

unam a dextris et alteram a sinistris eam quae a dextris erat on the left Boaz. uocauit Iachin et quae ad leuam Booz

4:1 fecit quoque altare aeneum uiginti cubitorum longitu- 4:1 Moreover he made an altar of brass, twenty cubits the

length thereof, and twenty cubits the breadth thereof, and ten cubits the height thereof.

dinis et uiginti cubitorum latitudinis et decem cubitorum altitudinis 4:2 mare etiam fusile decem cubitis a labio usque 4:2 Also he made a molten sea of ten cubits from brim to

brim, round in compass, and five cubits the height thereof; and a line of thirty cubits did compass it round about.

ad labium rotundum per circuitum quinque cubitos habebat altitudinis et funiculus triginta cubitorum ambiebat gyrum eius 4:3 similitudo quoque boum erat subter illud et decem 4:3 And under it was the similitude of oxen, which did

compass it round about: ten in a cubit, compassing the sea round about. Two rows of oxen were cast, when it was

cubitis quaedam extrinsecus celaturae quasi duobus uersibus cast. aluum maris circuibant boues autem erant fusiles 4:4 et ip- 4:4 It stood upon twelve oxen, three looking toward the

north, and three looking toward the west, and three looking toward the south, and three looking toward the east: and the sea was set above upon them, and all their hinder parts were inward.

sum mare super duodecim boues inpositum erat quorum tres respiciebant aquilonem et alii tres occidentem porro tres alii meridiem et tres qui reliqui erant orientem mare habentes superpositum posteriora autem boum erant intrinsecus sub mari 4:5 porro uastitas eius habebat mensuram palmi et la- 4:5 And the thickness of it was an handbreadth, and the

brim of it like the work of the brim of a cup, with flowers of lilies; and it received and held three thousand baths.

bium illius erat quasi labium calicis uel repandi lilii capiebatque mensurae tria milia metretas 4:6 fecit quoque con- 4:6 He made also ten lavers, and put five on the right hand,

and five on the left, to wash in them: such things as they offered for the burnt offering they washed in them; but the

cas decem et posuit quinque a dextris et quinque a sinistris sea was for the priests to wash in. ut lauarent in eis omnia quae in holocaustum oblaturi erant porro in mari sacerdotes lauabantur 4:7 fecit autem et can- 4:7 And he made ten candlesticks of gold according to

their form, and set them in the temple, five on the right hand, and five on the left.

delabra aurea decem secundum speciem qua iussa erant fieri et posuit ea in templo quinque a dextris et quinque a sinistris 4:8 necnon et mensas decem posuitque eas in templo quin- 4:8 He made also ten tables, and placed them in the tem-

ple, five on the right side, and five on the left. And he made an hundred basons of gold.

que a dextris et quinque a sinistris fialas quoque aureas centum 4:9 fecit etiam atrium sacerdotum et basilicam gran- 4:9 Furthermore he made the court of the priests, and the

great court, and doors for the court, and overlaid the doors of them with brass.

dem et ostia in basilica quae texit aere 4:10 porro mare po-

4:10 And he set the sea on the right side of the east end, over against the south.

suit in latere dextro contra orientem ad meridiem 4:11 fecit 4:11 And Huram made the pots, and the shovels, and the

basons. And Huram finished the work that he was to make for king Solomon for the house of God;

autem Hiram lebetas quoque et creagras et fialas et conpleuit omne opus regis in domo Dei 4:12 hoc est columnas duas 4:12 To wit, the two pillars, and the pommels, and the

chapiters which were on the top of the two pillars, and the two wreaths to cover the two pommels of the chapiters

et epistylia et capita et quasi quaedam retiacula quae capita which were on the top of the pillars; tegerent super epistylia 4:13 malagranata quoque quadrin- 4:13 And four hundred pomegranates on the two wreaths;

two rows of pomegranates on each wreath, to cover the two pommels of the chapiters which were upon the pillars.

Biblia Sacra Vulgata

genta et retiacula duo ita ut bini ordines malagranatorum singulis retiaculis iungerentur quae protegerent epistylia et ca-

4:14 He made also bases, and lavers made he upon the pita columnarum 4:14 bases etiam fecit et concas quas su-

bases;

4:15 One sea, and twelve oxen under it. perposuit basibus 4:15 mare unum bouesque duodecim sub 4:16 The pots also, and the shovels, and the fleshhooks, mari 4:16 et lebetas et creagras et fialas omnia uasa fecit

and all their instruments, did Huram his father make to king Solomon for the house of the LORD of bright brass.

Salomoni Hiram pater eius in domo Domini ex aere mun-

4:17 In the plain of Jordan did the king cast them, in the dissimo 4:17 in regione Iordanis fudit ea rex in argillosa

clay ground between Succoth and Zeredathah.

4:18 Thus Solomon made all these vessels in great abun- terra inter Socchoth et Saredatha 4:18 erat autem multitudo

dance: for the weight of the brass could not be found out.

4:19 And Solomon made all the vessels that were for the uasorum innumerabilis ita ut ignoraretur pondus aeris 4:19

house of God, the golden altar also, and the tables whereon the shewbread was set;

fecitque Salomon omnia uasa domus Dei et altare aureum

4:20 Moreover the candlesticks with their lamps, that they et mensas et super eas panes propositionis 4:20 candelabra should burn after the manner before the oracle, of pure gold;

quoque cum lucernis suis ut lucerent ante oraculum iuxta

4:21 And the flowers, and the lamps, and the tongs, made ritum ex auro purissimo 4:21 et florentia quaedam et lu-

he of gold, and that perfect gold;

cernas et forcipes aureos omnia de auro mundissimo facta

4:22 And the snuffers, and the basons, and the spoons, and sunt 4:22 thymiamateria quoque et turibula et fialas et mor-

the censers, of pure gold: and the entry of the house, the inner doors thereof for the most holy place, and the doors

of the house of the temple, were of gold. tariola ex auro purissimo et ostia celauit templi interioris id est in sancto sanctorum et ostia templi forinsecus aurea sicque conpletum est omne opus quod fecit Salomon in domo Domini

5:1 Thus all the work that Solomon made for the house of 5:1 intulit igitur Salomon omnia quae uouerat Dauid pater

the LORD was finished: and Solomon brought in all the things that David his father had dedicated; and the silver, and the gold, and all the instruments, put he among the treasures of the house of God.

suus argentum et aurum et uniuersa uasa posuit in thesauris

5:2 Then Solomon assembled the elders of Israel, and all domus Dei 5:2 post quae congregauit maiores natu Israhel

the heads of the tribes, the chief of the fathers of the children of Israel, unto Jerusalem, to bring up the ark of the covenant of the LORD out of the city of David, which is Zion.

et cunctos principes tribuum et capita familiarum de filiis Israhel in Hierusalem ut adducerent arcam foederis Domini

5:3 Wherefore all the men of Israel assembled themselves de ciuitate Dauid quae est Sion 5:3 uenerunt igitur ad reunto the king in the feast which was in the seventh month. 5:4 And all the elders of Israel came; and the Levites took gem omnes uiri Israhel in die sollemni mensis septimi 5:4 up the ark.

cumque uenissent cuncti seniorum Israhel portauerunt Leui-

5:5 And they brought up the ark, and the tabernacle of tae arcam 5:5 et intulerunt eam et omnem paraturam taber-

the congregation, and all the holy vessels that were in the tabernacle, these did the priests and the Levites bring up.

naculi porro uasa sanctuarii quae erant in tabernaculo por-

5:6 Also king Solomon, and all the congregation of Israel tauerunt sacerdotes cum Leuitis 5:6 rex autem Salomon et that were assembled unto him before the ark, sacrificed sheep and oxen, which could not be told nor numbered for

multitude. uniuersus coetus Israhel et omnes qui fuerant congregati ante

arcam immolabant arietes et boues absque ullo numero tanta

5:7 And the priests brought in the ark of the covenant of enim erat multitudo uictimarum 5:7 et intulerunt sacerdo-

the LORD unto his place, to the oracle of the house, into the most holy place, even under the wings of the cherubims:

Liber II Paralipomenon

tes arcam foederis Domini in locum suum id est ad oraculum templi in sancta sanctorum subter alas cherubin 5:8 ita ut 5:8 For the cherubims spread forth their wings over the

place of the ark, and the cherubims covered the ark and the staves thereof above.

cherubin expanderent alas suas super locum in quo posita erat arca et ipsam arcam tegerent cum uectibus eius 5:9 5:9 And they drew out the staves of the ark, that the ends

of the staves were seen from the ark before the oracle; but they were not seen without. And there it is unto this day.

uectium autem quibus portabatur arca quia paululum longiores erant capita parebant ante oraculum si uero quis paululum fuisset extrinsecus eos uidere non poterat fuit itaque arca ibi usque in praesentem diem 5:10 nihilque erat al- 5:10 There was nothing in the ark save the two tables

which Moses put therein at Horeb, when the LORD made a covenant with the children of Israel, when they came out

iud in arca nisi duae tabulae quas posuerat Moses in Ho- of Egypt. reb quando legem dedit Dominus filiis Israhel egredientibus ex Aegypto 5:11 egressis autem sacerdotibus de sanctuario 5:11 And it came to pass, when the priests were come out

of the holy place: (for all the priests that were present were sanctified, and did not then wait by course:

omnes enim sacerdotes qui ibi potuerant inueniri sanctificati sunt nec adhuc illo tempore uices et ministeriorum ordo inter eos diuisus erat 5:12 tam Leuitae quam cantores id est 5:12 Also the Levites which were the singers, all of them

of Asaph, of Heman, of Jeduthun, with their sons and their brethren, being arrayed in white linen, having cymbals and psalteries and harps, stood at the east end of the altar, and with them an hundred and twenty priests sounding with trumpets:)

et qui sub Asaph erant et qui sub Heman et qui sub Idithun filii et fratres eorum uestiti byssinis cymbalis et psalteriis et citharis concrepabant stantes ad orientalem plagam altaris cumque eis sacerdotes centum uiginti canentes tubis 5:13 igitur cunctis pariter et tubis et uoce et cymbalis et or- 5:13 It came even to pass, as the trumpeters and singers

were as one, to make one sound to be heard in praising and thanking the LORD; and when they lifted up their voice with the trumpets and cymbals and instruments of musick, and praised the LORD, saying, For he is good; for his mercy endureth for ever: that then the house was filled with a cloud, even the house of the LORD;

ganis et diuersi generis musicorum concinentibus et uocem in sublime tollentibus longe sonitus audiebatur ita ut cum Dominum laudare coepissent et dicere confitemini Domino quoniam bonus quoniam in aeternum misericordia eius im-

pleretur domus Domini nube 5:14 nec possent sacerdotes 5:14 So that the priests could not stand to minister by reason of the cloud: for the glory of the LORD had filled the house of God.

stare et ministrare propter caliginem conpleuerat enim gloria Domini domum Dei

6:1 tunc Salomon ait Dominus pollicitus est ut habitaret 6:1 Then said Solomon, The LORD hath said that he

would dwell in the thick darkness.

in caligine 6:2 ego autem aedificaui domum nomini eius 6:2 But I have built an house of habitation for thee, and a

place for thy dwelling for ever.

ut habitaret ibi in perpetuum 6:3 et conuertit faciem suam 6:3 And the king turned his face, and blessed the whole

congregation of Israel: and all the congregation of Israel stood.

et benedixit uniuersae multitudini Israhel nam omnis turba stabat intenta et ait 6:4 benedictus Dominus Deus Israhel 6:4 And he said, Blessed be the LORD God of Israel, who

hath with his hands fulfilled that which he spake with his mouth to my father David, saying,

qui quod locutus est Dauid patri meo opere conpleuit dicens

6:5 a die qua eduxi populum meum de terra Aegypti non 6:5 Since the day that I brought forth my people out of the land of Egypt I chose no city among all the tribes of Israel to build an house in, that my name might be there; neither chose I any man to be a ruler over my people Israel:

Biblia Sacra Vulgata

elegi ciuitatem de cunctis tribubus Israhel ut aedificaretur in ea domus nomini meo neque elegi quemquam alium uirum

6:6 But I have chosen Jerusalem, that my name might be ut esset dux in populo meo Israhel 6:6 sed elegi Hierusalem

there; and have chosen David to be over my people Israel.

ut sit nomen meum in ea et elegi Dauid ut constituerem eum

6:7 Now it was in the heart of David my father to build an super populum meum Israhel 6:7 cumque fuisset uoluntatis

house for the name of the LORD God of Israel.

Dauid patris mei ut aedificaret domum nomini Domini Dei

6:8 But the LORD said to David my father, Forasmuch as Israhel 6:8 dixit Dominus ad eum quia haec fuit uoluntas

it was in thine heart to build an house for my name, thou didst well in that it was in thine heart:

tua ut aedificares domum nomini meo bene quidem fecisti

6:9 Notwithstanding thou shalt not build the house; but thy habere huiuscemodi uoluntatem 6:9 sed non tu aedificabis son which shall come forth out of thy loins, he shall build the house for my name.

domum uerum filius tuus qui egredietur de lumbis tuis ipse

6:10 The LORD therefore hath performed his word that he aedificabit domum nomini meo 6:10 conpleuit ergo Domihath spoken: for I am risen up in the room of David my father, and am set on the throne of Israel, as the LORD promised, and have built the house for the name of the LORD God of Israel.

nus sermonem suum quem locutus fuerat et ego surrexi pro Dauid patre meo et sedi super thronum Israhel sicut locutus est Dominus et aedificaui domum nomini Domini Dei

6:11 And in it have I put the ark, wherein is the covenant Israhel 6:11 et posui in ea arcam in qua est pactum Domini

of the LORD, that he made with the children of Israel.

6:12 And he stood before the altar of the LORD in the quod pepigit cum filiis Israhel 6:12 stetit ergo coram altare

presence of all the congregation of Israel, and spread forth his hands:

Domini ex aduerso uniuersae multitudinis Israhel et extendit

6:13 For Solomon had made a brasen scaffold of five cu- manus suas 6:13 siquidem fecerat Salomon basem aeneam

bits long, and five cubits broad, and three cubits high, and had set it in the midst of the court: and upon it he stood, and kneeled down upon his knees before all the congregation of Israel, and spread forth his hands toward heaven.

et posuerat eam in medio basilicae habentem quinque cubitos longitudinis et quinque cubitos latitudinis et tres cubitos in altum stetitque super eam et deinceps flexis genibus contra uniuersam multitudinem Israhel et palmis in caelum leua-

6:14 And said, O LORD God of Israel, there is no God tis 6:14 ait Domine Deus Israhel non est similis tui Deus in

like thee in the heaven, nor in the earth; which keepest covenant, and shewest mercy unto thy servants, that walk

before thee with all their hearts: caelo et in terra qui custodis pactum et misericordiam cum 6:15 Thou which hast kept with thy servant David my fa- seruis tuis qui ambulant coram te in toto corde suo 6:15

ther that which thou hast promised him; and spakest with thy mouth, and hast fulfilled it with thine hand, as it is this

day. qui praestitisti seruo tuo Dauid patri meo quaecumque lo-

cutus fueras ei et quae ore promiseras opere conplesti sicut

6:16 Now therefore, O LORD God of Israel, keep with et praesens tempus probat 6:16 nunc ergo Domine Deus Is-

thy servant David my father that which thou hast promised him, saying, There shall not fail thee a man in my sight to sit upon the throne of Israel; yet so that thy children take heed to their way to walk in my law, as thou hast walked before me.

rahel imple seruo tuo patri meo Dauid quaecumque locutus es dicens non deficiet ex te uir coram me qui sedeat super thronum Israhel ita tamen si custodierint filii tui uias suas et ambulauerint in lege mea sicut et tu ambulasti coram me

6:17 Now then, O LORD God of Israel, let thy word be 6:17 et nunc Domine Deus Israhel firmetur sermo tuus quem

verified, which thou hast spoken unto thy servant David.

Liber II Paralipomenon

locutus es seruo tuo Dauid 6:18 ergone credibile est ut ha- 6:18 But will God in very deed dwell with men on the

earth? behold, heaven and the heaven of heavens cannot contain thee; how much less this house which I have built!

bitet Deus cum hominibus super terram si caelum et caeli caelorum non te capiunt quanto magis domus ista quam aedificaui 6:19 sed ad hoc tantum facta est ut respicias ora- 6:19 Have respect therefore to the prayer of thy servant,

and to his supplication, O LORD my God, to hearken unto the cry and the prayer which thy servant prayeth before

tionem serui tui et obsecrationem eius Domine Deus meus thee: audias et preces quas fundit famulus tuus coram te 6:20 6:20 That thine eyes may be open upon this house day

and night, upon the place whereof thou hast said that thou wouldest put thy name there; to hearken unto the prayer

ut aperias oculos tuos super domum istam diebus et nocti- which thy servant prayeth toward this place. bus super locum in quo pollicitus es ut inuocaretur nomen

tuum 6:21 et exaudires orationem quam seruus tuus orat in 6:21 Hearken therefore unto the supplications of thy servant, and of thy people Israel, which they shall make toward this place: hear thou from thy dwelling place, even

eo exaudi preces famuli tui et populi tui Israhel quicumque from heaven; and when thou hearest, forgive. orauerit in loco isto et exaudi de habitaculo tuo id est de caelis et propitiare 6:22 si peccauerit quispiam in proximum 6:22 If a man sin against his neighbour, and an oath be laid

upon him to make him swear, and the oath come before thine altar in this house;

suum et iurare contra eum paratus uenerit seque maledicto constrinxerit coram altari in domo ista 6:23 tu audies de 6:23 Then hear thou from heaven, and do, and judge thy

servants, by requiting the wicked, by recompensing his way upon his own head; and by justifying the righteous,

caelo et facies iudicium seruorum tuorum ita ut reddas in- by giving him according to his righteousness. iquo uiam suam in caput proprium et ulciscaris iustum retribuens ei secundum iustitiam suam 6:24 si superatus fuerit 6:24 And if thy people Israel be put to the worse before the

enemy, because they have sinned against thee; and shall return and confess thy name, and pray and make supplica-

populus tuus Israhel ab inimicis peccabunt enim tibi et con- tion before thee in this house; uersi egerint paenitentiam et obsecrauerint nomen tuum et fuerint deprecati in loco isto 6:25 tu exaudi de caelo et pro- 6:25 Then hear thou from the heavens, and forgive the sin

of thy people Israel, and bring them again unto the land which thou gavest to them and to their fathers.

pitiare peccato populi tui Israhel et reduc eos in terram quam dedisti eis et patribus eorum 6:26 si clauso caelo pluuia non 6:26 When the heaven is shut up, and there is no rain, be-

cause they have sinned against thee; yet if they pray toward this place, and confess thy name, and turn from their

fluxerit propter peccata populi et deprecati te fuerint in loco sin, when thou dost afflict them; isto et confessi nomini tuo et conuersi a peccatis suis cum eos adflixeris 6:27 exaudi de caelo Domine et dimitte pec- 6:27 Then hear thou from heaven, and forgive the sin of

thy servants, and of thy people Israel, when thou hast taught them the good way, wherein they should walk; and send rain upon thy land, which thou hast given unto thy people for an inheritance.

cata seruis tuis et populi tui Israhel et doce eos uiam bonam per quam ingrediantur et da pluuiam terrae quam dedisti po-

pulo tuo ad possidendum 6:28 fames si orta fuerit in terra 6:28 If there be dearth in the land, if there be pestilence, if there be blasting, or mildew, locusts, or caterpillers; if their enemies besiege them in the cities of their land; what-

et pestilentia erugo et aurugo et lucusta et brucus et hos- soever sore or whatsoever sickness there be: tes uastatis regionibus portas obsederint ciuitatis omnisque plaga et infirmitas presserit 6:29 si quis de populo tuo Is- 6:29 Then what prayer or what supplication soever shall be

made of any man, or of all thy people Israel, when every one shall know his own sore and his own grief, and shall

rahel fuerit deprecatus cognoscens plagam et infirmitatem spread forth his hands in this house: suam et expanderit manus suas in domo hac 6:30 tu exaudi 6:30 Then hear thou from heaven thy dwelling place, and

forgive, and render unto every man according unto all his ways, whose heart thou knowest; (for thou only knowest the hearts of the children of men:)

Biblia Sacra Vulgata

de caelo de sublimi scilicet habitaculo tuo et propitiare et redde unicuique secundum uias suas quas nosti eum habere in corde suo tu enim solus nosti corda filiorum hominum

6:31 That they may fear thee, to walk in thy ways, so long 6:31 ut timeant te et ambulent in uiis tuis cunctis diebus qui-

as they live in the land which thou gavest unto our fathers.

bus uiuunt super faciem terrae quam dedisti patribus nostris

6:32 Moreover concerning the stranger, which is not of thy 6:32 externum quoque qui non est de populo tuo Israhel si people Israel, but is come from a far country for thy great name’s sake, and thy mighty hand, and thy stretched out

arm; if they come and pray in this house; uenerit de terra longinqua propter nomen tuum magnum et

propter manum tuam robustam et brachium tuum extentum

6:33 Then hear thou from the heavens, even from thy et adorauerit in loco isto 6:33 tu exaudies de caelo firmis-

dwelling place, and do according to all that the stranger calleth to thee for; that all people of the earth may know thy name, and fear thee, as doth thy people Israel, and may know that this house which I have built is called by thy name.

simo habitaculo tuo et facies cuncta pro quibus inuocauerit te ille peregrinus ut sciant omnes populi terrae nomen tuum et timeant te sicut populus tuus Israhel et cognoscant quia nomen tuum inuocatum est super domum hanc quam aedi-

6:34 If thy people go out to war against their enemies by ficaui 6:34 si egressus fuerit populus tuus ad bellum contra

the way that thou shalt send them, and they pray unto thee toward this city which thou hast chosen, and the house

which I have built for thy name; aduersarios suos per uiam in qua miseris eos adorabunt te

contra uiam in qua ciuitas haec est quam elegisti et domus

6:35 Then hear thou from the heavens their prayer and quam aedificaui nomini tuo 6:35 ut exaudias de caelo pre-

their supplication, and maintain their cause.

6:36 If they sin against thee, (for there is no man which ces eorum et obsecrationem et ulciscaris 6:36 si autem et

sinneth not,) and thou be angry with them, and deliver them over before their enemies, and they carry them away

captives unto a land far off or near; peccauerint tibi neque enim est homo qui non peccet et ira-

tus fueris eis et tradideris hostibus et captiuos eos duxerint

6:37 Yet if they bethink themselves in the land whither in terram longinquam uel certe quae iuxta est 6:37 et con-

they are carried captive, and turn and pray unto thee in the land of their captivity, saying, We have sinned, we have

done amiss, and have dealt wickedly; uersi corde suo in terra ad quam captiui ducti fuerant egerint

paenitentiam et deprecati te fuerint in terra captiuitatis suae

6:38 If they return to thee with all their heart and with all dicentes peccauimus inique fecimus iniuste egimus 6:38 et

their soul in the land of their captivity, whither they have carried them captives, and pray toward their land, which thou gavest unto their fathers, and toward the city which thou hast chosen, and toward the house which I have built for thy name:

reuersi fuerint ad te in toto corde suo et in tota anima sua in terra captiuitatis suae ad quam ducti sunt adorabunt te contra uiam terrae suae quam dedisti patribus eorum et urbis quam

6:39 Then hear thou from the heavens, even from thy elegisti et domus quam aedificaui nomini tuo 6:39 ut exau-

dwelling place, their prayer and their supplications, and maintain their cause, and forgive thy people which have

sinned against thee. dias de caelo hoc est de firmo habitaculo tuo preces eorum

et facias iudicium et dimittas populo tuo quamuis peccatori

6:40 Now, my God, let, I beseech thee, thine eyes be open, 6:40 tu es enim Deus meus aperiantur quaeso oculi tui et au-

and let thine ears be attent unto the prayer that is made in this place.

res tuae intentae sint ad orationem quae fit in loco isto 6:41

6:41 Now therefore arise, O LORD God, into thy resting place, thou, and the ark of thy strength: let thy priests, O LORD God, be clothed with salvation, and let thy saints rejoice in goodness.

nunc igitur consurge Domine Deus in requiem tuam tu et

Liber II Paralipomenon

arca fortitudinis tuae sacerdotes tui Domine Deus induantur salute et sancti tui laetentur in bonis 6:42 Domine Deus ne 6:42 O LORD God, turn not away the face of thine

anointed: remember the mercies of David thy servant.

auerseris faciem christi tui memento misericordiarum Dauid serui tui

7:1 cumque conplesset Salomon fundens preces ignis de- 7:1 Now when Solomon had made an end of praying, the

fire came down from heaven, and consumed the burnt offering and the sacrifices; and the glory of the LORD filled

scendit de caelo et deuorauit holocausta et uictimas et maies- the house. tas Domini impleuit domum 7:2 nec poterant sacerdotes in- 7:2 And the priests could not enter into the house of the

LORD, because the glory of the LORD had filled the LORD’s house.

gredi templum Domini eo quod implesset maiestas Domini templum Domini 7:3 sed et omnes filii Israhel uidebant de- 7:3 And when all the children of Israel saw how the fire

came down, and the glory of the LORD upon the house, they bowed themselves with their faces to the ground upon the pavement, and worshipped, and praised the LORD, saying, For he is good; for his mercy endureth for ever.

scendentem ignem et gloriam Domini super domum et corruentes proni in terram super pauimentum stratum lapide adorauerunt et laudauerunt Dominum quoniam bonus quoniam in aeternum misericordia eius 7:4 rex autem et omnis popu- 7:4 Then the king and all the people offered sacrifices be-

fore the LORD.

lus immolabant uictimas coram Domino 7:5 mactauit igitur 7:5 And king Solomon offered a sacrifice of twenty and

two thousand oxen, and an hundred and twenty thousand sheep: so the king and all the people dedicated the house

rex Salomon hostias boum uiginti duo milia arietum centum of God. uiginti milia et dedicauit domum Dei rex et uniuersus populus 7:6 sacerdotes autem stabant in officiis suis et Leuitae 7:6 And the priests waited on their offices: the Levites also

with instruments of musick of the LORD, which David the king had made to praise the LORD, because his mercy endureth for ever, when David praised by their ministry; and the priests sounded trumpets before them, and all Israel stood.

in organis carminum Domini quae fecit Dauid rex ad laudandum Dominum quoniam in aeternum misericordia eius hymnos Dauid canentes per manus suas porro sacerdotes canebant tubis ante eos cunctusque Israhel stabat 7:7 sancti- 7:7 Moreover Solomon hallowed the middle of the court

that was before the house of the LORD: for there he offered burnt offerings, and the fat of the peace offerings, because the brasen altar which Solomon had made was not able to receive the burnt offerings, and the meat offerings, and the fat.

ficauit quoque Salomon medium atrii ante templum Domini obtulerat enim ibi holocausta et adipes pacificorum quia altare aeneum quod fecerat non poterat sustinere holocausta et sacrificia et adipes 7:8 fecit ergo Salomon sollemnita- 7:8 Also at the same time Solomon kept the feast seven

days, and all Israel with him, a very great congregation, from the entering in of Hamath unto the river of Egypt.

tem in tempore illo septem diebus et omnis Israhel cum eo ecclesia magna ualde ab introitu Emath usque ad torrentem Aegypti 7:9 fecitque die octauo collectam eo quod dedicas- 7:9 And in the eighth day they made a solemn assembly:

for they kept the dedication of the altar seven days, and the feast seven days.

set altare septem diebus et sollemnitatem celebrasset diebus septem 7:10 igitur in die uicesimo tertio mensis septimi di- 7:10 And on the three and twentieth day of the seventh

month he sent the people away into their tents, glad and merry in heart for the goodness that the LORD had shewed

misit populos ad tabernacula sua laetantes atque gaudentes unto David, and to Solomon, and to Israel his people. super bono quod fecerat Dominus Dauid et Salomoni et Israhel populo suo 7:11 conpleuitque Salomon domum Do- 7:11 Thus Solomon finished the house of the LORD, and

the king’s house: and all that came into Solomon’s heart to make in the house of the LORD, and in his own house, he prosperously effected.

Biblia Sacra Vulgata

mini et domum regis et omnia quae disposuerat in corde suo ut faceret in domo Domini et in domo sua et prosperatus

7:12 And the LORD appeared to Solomon by night, and est 7:12 apparuit autem ei Dominus nocte et ait audiui ora-

said unto him, I have heard thy prayer, and have chosen this place to myself for an house of sacrifice.

tionem tuam et elegi locum istum mihi in domum sacrificii

7:13 If I shut up heaven that there be no rain, or if I com- 7:13 si clausero caelum et pluuia non fluxerit et mandauero et

mand the locusts to devour the land, or if I send pestilence among my people;

praecepero lucustae ut deuoret terram et misero pestilentiam

7:14 If my people, which are called by my name, shall in populum meum 7:14 conuersus autem populus meus su-

humble themselves, and pray, and seek my face, and turn from their wicked ways; then will I hear from heaven, and

will forgive their sin, and will heal their land. per quos inuocatum est nomen meum deprecatus me fuerit et exquisierit faciem meam et egerit paenitentiam a uiis suis pessimis et ego exaudiam de caelo et propitius ero peccatis

7:15 Now mine eyes shall be open, and mine ears attent eorum et sanabo terram eorum 7:15 oculi quoque mei erunt

unto the prayer that is made in this place.

aperti et aures meae erectae ad orationem eius qui in loco

7:16 For now have I chosen and sanctified this house, that isto orauerit 7:16 elegi enim et sanctificaui locum istum ut

my name may be there for ever: and mine eyes and mine heart shall be there perpetually.

sit nomen meum ibi in sempiternum et permaneant oculi mei

7:17 And as for thee, if thou wilt walk before me, as David et cor meum ibi cunctis diebus 7:17 tu quoque si ambulathy father walked, and do according to all that I have commanded thee, and shalt observe my statutes and my judg-

ments; ueris coram me sicut ambulauit Dauid pater tuus et feceris

iuxta omnia quae praecepi tibi et iustitias meas iudiciaque

7:18 Then will I stablish the throne of thy kingdom, ac- seruaueris 7:18 suscitabo thronum regni tui sicut pollicitus

cording as I have covenanted with David thy father, saying, There shall not fail thee a man to be ruler in Israel.

sum Dauid patri tuo dicens non auferetur de stirpe tua uir

7:19 But if ye turn away, and forsake my statutes and my qui sit princeps in Israhel 7:19 si autem auersi fueritis et

commandments, which I have set before you, and shall go and serve other gods, and worship them;

dereliqueritis iustitias meas et praecepta mea quae proposui uobis et abeuntes seruieritis diis alienis et adoraueritis eos

7:20 Then will I pluck them up by the roots out of my 7:20 euellam uos de terra mea quam dedi uobis et domum

land which I have given them; and this house, which I have sanctified for my name, will I cast out of my sight, and will make it to be a proverb and a byword among all nations.

hanc quam sanctificaui nomini meo proiciam a facie mea et tradam eam in parabolam et in exemplum cunctis populis

7:21 And this house, which is high, shall be an astonish- 7:21 et domus ista erit in prouerbium uniuersis transeuntibus

ment to every one that passeth by it; so that he shall say, Why hath the LORD done thus unto this land, and unto

this house? et dicent stupentes quare fecit Dominus sic terrae huic et do-7:22 And it shall be answered, Because they forsook the mui huic 7:22 respondebuntque quia dereliquerunt Domi-

LORD God of their fathers, which brought them forth out of the land of Egypt, and laid hold on other gods, and worshipped them, and served them: therefore hath he brought all this evil upon them.

num Deum patrum suorum qui eduxit eos de terra Aegypti et adprehenderunt deos alienos et adorauerunt eos atque coluerunt idcirco uenerunt super eos uniuersa haec mala

8:1 And it came to pass at the end of twenty years, wherein 8:1 expletis autem uiginti annis postquam aedificauit Sa-

Solomon had built the house of the LORD, and his own house,

lomon domum Domini et domum suam 8:2 ciuitates quas

8:2 That the cities which Huram had restored to Solomon, Solomon built them, and caused the children of Israel to dwell there.

Liber II Paralipomenon

dederat Hiram Salomoni aedificauit et habitare ibi fecit filios Israhel 8:3 abiit quoque in Emath Suba et obtinuit eam 8:3 And Solomon went to Hamathzobah, and prevailed

against it.

8:4 et aedificauit Palmyram in deserto et alias ciuitates muni- 8:4 And he built Tadmor in the wilderness, and all the store

cities, which he built in Hamath.

tissimas aedificauit in Emath 8:5 extruxitque Bethoron su- 8:5 Also he built Bethhoron the upper, and Bethhoron the

nether, fenced cities, with walls, gates, and bars;

periorem et Bethoron inferiorem ciuitates muratas habentes

portas et uectes et seras 8:6 Baalath etiam et omnes urbes 8:6 And Baalath, and all the store cities that Solomon had, and all the chariot cities, and the cities of the horsemen, and all that Solomon desired to build in Jerusalem, and in

firmissimas quae fuerunt Salomonis cunctasque urbes quad- Lebanon, and throughout all the land of his dominion. rigarum et urbes equitum omnia quae uoluit Salomon atque disposuit aedificauit in Hierusalem et in Libano et in uni-

uersa terra potestatis suae 8:7 omnem populum qui dere- 8:7 As for all the people that were left of the Hittites, and the Amorites, and the Perizzites, and the Hivites, and the Jebusites, which were not of Israel,

lictus fuerat de Hettheis et Amorreis et Ferezeis et Eueis et Iebuseis qui non erant de stirpe Israhel 8:8 de filiis eorum 8:8 But of their children, who were left after them in the

land, whom the children of Israel consumed not, them did Solomon make to pay tribute until this day.

et de posteris quos non interfecerant filii Israhel subiugauit Salomon in tributarios usque in diem hanc 8:9 porro de fi- 8:9 But of the children of Israel did Solomon make no

servants for his work; but they were men of war, and chief of his captains, and captains of his chariots and horsemen.

liis Israhel non posuit ut seruirent operibus regis ipsi enim erant uiri bellatores et duces primi et principes quadrigarum et equitum eius 8:10 omnes autem principes exercitus regis 8:10 And these were the chief of king Solomon’s officers,

even two hundred and fifty, that bare rule over the people.

Salomonis fuerunt ducenti quinquaginta qui erudiebant populum 8:11 filiam uero Pharaonis transtulit de ciuitate Da- 8:11 And Solomon brought up the daughter of Pharaoh out

of the city of David unto the house that he had built for her: for he said, My wife shall not dwell in the house of David king of Israel, because the places are holy, whereunto the ark of the LORD hath come.

uid in domum quam aedificauerat ei dixit enim non habitabit uxor mea in domo Dauid regis Israhel eo quod sanctificata sit quia ingressa est eam arca Domini 8:12 tunc ob- 8:12 Then Solomon offered burnt offerings unto the

LORD on the altar of the LORD, which he had built before the porch,

tulit Salomon holocausta Domino super altare Domini quod extruxerat ante porticum 8:13 ut per singulos dies offerre- 8:13 Even after a certain rate every day, offering according

to the commandment of Moses, on the sabbaths, and on the new moons, and on the solemn feasts, three times in the year, even in the feast of unleavened bread, and in the feast of weeks, and in the feast of tabernacles.

tur in eo iuxta praeceptum Mosi in sabbatis et in kalendis et in festis diebus ter per annum id est in sollemnitate azymorum et in sollemnitate ebdomadarum et in sollemnitate tabernaculorum 8:14 et constituit iuxta dispositionem Da- 8:14 And he appointed, according to the order of David

his father, the courses of the priests to their service, and the Levites to their charges, to praise and minister before the priests, as the duty of every day required: the porters also by their courses at every gate: for so had David the man of God commanded.

uid patris sui officia sacerdotum in ministeriis suis et Leuitas in ordine suo ut laudarent et ministrarent coram sacerdotibus iuxta ritum uniuscuiusque diei et ianitores in diuisionibus suis per portam et portam sic enim praeceperat Dauid homo Dei 8:15 nec praetergressi sunt de mandatis regis 8:15 And they departed not from the commandment of the

king unto the priests and Levites concerning any matter, or concerning the treasures.

Biblia Sacra Vulgata

tam sacerdotes quam Leuitae ex omnibus quae praeceperat

8:16 Now all the work of Solomon was prepared unto the et in custodiis thesaurorum 8:16 omnes inpensas praepa-

day of the foundation of the house of the LORD, and until it was finished. So the house of the LORD was perfected.

ratas habuit Salomon ex eo die quo fundauit domum Do-

8:17 Then went Solomon to Eziongeber, and to Eloth, at mini usque in diem quo perfecit eam 8:17 tunc abiit Sathe sea side in the land of Edom.

lomon in Hesiongaber et in Ahilath ad oram maris Rubri

8:18 And Huram sent him by the hands of his servants quae est in terra Edom 8:18 misit autem ei Hiram per ma-

ships, and servants that had knowledge of the sea; and they went with the servants of Solomon to Ophir, and took thence four hundred and fifty talents of gold, and brought them to king Solomon.

num seruorum suorum naues et nautas gnaros maris et abierunt cum seruis Salomonis in Ophir tuleruntque inde quadringenta quinquaginta talenta auri et adtulerunt ad regem Salomonem

9:1 And when the queen of Sheba heard of the fame of 9:1 regina quoque Saba cum audisset famam Salomonis

Solomon, she came to prove Solomon with hard questions at Jerusalem, with a very great company, and camels that bare spices, and gold in abundance, and precious stones: and when she was come to Solomon, she communed with him of all that was in her heart.

uenit ut temptaret eum enigmatibus in Hierusalem cum magnis opibus et camelis qui portabant aromata et auri plurimum gemmasque pretiosas cumque uenisset ad Salomonem

9:2 And Solomon told her all her questions: and there was locuta est ei quaecumque erant in corde suo 9:2 et exposuit

nothing hid from Solomon which he told her not.

ei Salomon omnia quae proposuerat nec quicquam fuit quod

9:3 And when the queen of Sheba had seen the wisdom of ei non perspicuum fecerit 9:3 quod postquam uidit sapien-

Solomon, and the house that he had built,

9:4 And the meat of his table, and the sitting of his ser- tiam scilicet Salomonis et domum quam aedificauerat 9:4

vants, and the attendance of his ministers, and their apparel; his cupbearers also, and their apparel; and his ascent by which he went up into the house of the LORD; there was no more spirit in her.

necnon cibaria mensae eius et habitacula seruorum et officia ministrorum eius et uestimenta eorum pincernas quoque et uestes eorum et uictimas quas immolabat in domo Do-

9:5 And she said to the king, It was a true report which I mini non erat prae stupore ultra in ea spiritus 9:5 dixitque

heard in mine own land of thine acts, and of thy wisdom:

ad regem uerus sermo quem audieram in terra mea de uirtu-

9:6 Howbeit I believed not their words, until I came, and tibus et sapientia tua 9:6 non credebam narrantibus donec

mine eyes had seen it: and, behold, the one half of the greatness of thy wisdom was not told me: for thou ex-

ceedest the fame that I heard. ipsa uenissem et uidissent oculi mei et probassem uix me-

dietatem mihi sapientiae tuae fuisse narratam uicisti famam

9:7 Happy are thy men, and happy are these thy servants, uirtutibus tuis 9:7 beati uiri tui et beati serui tui hii qui ad-

which stand continually before thee, and hear thy wisdom.

sistunt coram te in omni tempore et audiunt sapientiam tuam

9:8 Blessed be the LORD thy God, which delighted in thee 9:8 sit Dominus Deus tuus benedictus qui uoluit te ordinare

to set thee on his throne, to be king for the LORD thy God: because thy God loved Israel, to establish them for ever, therefore made he thee king over them, to do judgment and justice.

super thronum suum regem Domini Dei tui quia diligit Deus Israhel et uult seruare eum in aeternum idcirco posuit te

9:9 And she gave the king an hundred and twenty talents of super eum regem ut facias iudicia atque iustitiam 9:9 de-

gold, and of spices great abundance, and precious stones: neither was there any such spice as the queen of Sheba

gave king Solomon. dit autem regi centum uiginti talenta auri et aromata multa

Liber II Paralipomenon

nimis et gemmas pretiosissimas non fuerunt aromata talia ut haec quae dedit regina Saba regi Salomoni 9:10 sed et serui 9:10 And the servants also of Huram, and the servants of

Solomon, which brought gold from Ophir, brought algum trees and precious stones.

Hiram cum seruis Salomonis adtulerunt aurum de Ophir et ligna thyina et gemmas pretiosissimas 9:11 de quibus fe- 9:11 And the king made of the algum trees terraces to the

house of the LORD, and to the king’s palace, and harps and psalteries for singers: and there were none such seen

cit rex de lignis scilicet thyinis gradus in domo Domini et before in the land of Judah. in domo regia citharas quoque et psalteria cantoribus numquam uisa sunt in terra Iuda ligna talia 9:12 rex autem Salo- 9:12 And king Solomon gave to the queen of Sheba all her

desire, whatsoever she asked, beside that which she had brought unto the king. So she turned, and went away to

mon dedit reginae Saba cuncta quae uoluit et quae postulauit her own land, she and her servants. multo plura quam adtulerat ad eum quae reuersa abiit in terram suam cum seruis suis 9:13 erat autem pondus auri quod 9:13 Now the weight of gold that came to Solomon in

one year was six hundred and threescore and six talents of gold;

adferebatur Salomoni per annos singulos sescenta sexaginta sex talenta auri 9:14 excepta ea summa quam legati diuer- 9:14 Beside that which chapmen and merchants brought.

And all the kings of Arabia and governors of the country brought gold and silver to Solomon.

sarum gentium et negotiatores adferre consueuerant omnesque reges Arabiae et satrapae terrarum qui conportabant aurum et argentum Salomoni 9:15 fecit igitur rex Salomon 9:15 And king Solomon made two hundred targets of

beaten gold: six hundred shekels of beaten gold went to one target.

ducentas hastas aureas de summa sescentorum aureorum qui in hastis singulis expendebantur 9:16 trecenta quoque scuta 9:16 And three hundred shields made he of beaten gold:

three hundred shekels of gold went to one shield. And the king put them in the house of the forest of Lebanon.

aurea trecentorum aureorum quibus tegebantur scuta singula posuitque ea rex in armamentario quod erat consitum nemore 9:17 fecit quoque rex solium eburneum grande et ue- 9:17 Moreover the king made a great throne of ivory, and

overlaid it with pure gold.

stiuit illud auro mundissimo 9:18 sexque gradus quibus as- 9:18 And there were six steps to the throne, with a foot-

stool of gold, which were fastened to the throne, and stays on each side of the sitting place, and two lions standing by

cendebatur ad solium et scabillum aureum et brachiola duo the stays: altrinsecus et duos leones stantes iuxta brachiola 9:19 sed 9:19 And twelve lions stood there on the one side and on

the other upon the six steps. There was not the like made in any kingdom.

et alios duodecim leunculos stantes super sex gradus ex utraque parte non fuit tale solium in uniuersis regnis 9:20 om- 9:20 And all the drinking vessels of king Solomon were

of gold, and all the vessels of the house of the forest of Lebanon were of pure gold: none were of silver; it was

nia quoque uasa conuiuii regis erant aurea et uasa domus not any thing accounted of in the days of Solomon. saltus Libani ex auro purissimo argentum enim in diebus illis pro nihilo reputabatur 9:21 siquidem naues regis ibant 9:21 For the king’s ships went to Tarshish with the ser-

vants of Huram: every three years once came the ships of Tarshish bringing gold, and silver, ivory, and apes, and

in Tharsis cum seruis Hiram semel in annis tribus et defere- peacocks. bant inde aurum et argentum et ebur et simias et pauos 9:22 9:22 And king Solomon passed all the kings of the earth

in riches and wisdom.

magnificatus est igitur Salomon super omnes reges terrae diuitiis et gloria 9:23 omnesque reges terrarum desiderabant 9:23 And all the kings of the earth sought the presence

of Solomon, to hear his wisdom, that God had put in his heart.

faciem uidere Salomonis ut audirent sapientiam quam dede-

Biblia Sacra Vulgata

9:24 And they brought every man his present, vessels rat Deus in corde eius 9:24 et deferebant ei munera uasa

of silver, and vessels of gold, and raiment, harness, and spices, horses, and mules, a rate year by year.

argentea et aurea et uestes et arma et aromata equos et mu-

9:25 And Solomon had four thousand stalls for horses and los per singulos annos 9:25 habuit quoque Salomon quachariots, and twelve thousand horsemen; whom he bestowed in the chariot cities, and with the king at Jerusalem.

draginta milia equorum in stabulis et curruum equitumque duodecim milia constituitque eos in urbibus quadrigarum et

9:26 And he reigned over all the kings from the river even ubi erat rex in Hierusalem 9:26 exercuit etiam potestatem

unto the land of the Philistines, and to the border of Egypt.

super cunctos reges a fluuio Eufraten usque ad terram Philis-

9:27 And the king made silver in Jerusalem as stones, and thinorum id est usque ad terminos Aegypti 9:27 tantamque

cedar trees made he as the sycomore trees that are in the low plains in abundance.

copiam praebuit argenti in Hierusalem quasi lapidum et cedrorum tantam multitudinem uelut sycaminorum quae gig-

9:28 And they brought unto Solomon horses out of Egypt, nuntur in campestribus 9:28 adducebantur autem ei equi de

and out of all lands. 9:29 Now the rest of the acts of Solomon, first and last, are Aegypto cunctisque regionibus 9:29 reliqua uero operum they not written in the book of Nathan the prophet, and in the prophecy of Ahijah the Shilonite, and in the visions of

Iddo the seer against Jeroboam the son of Nebat? Salomonis priorum et nouissimorum scripta sunt in uerbis

Nathan prophetae et in libris Ahiae Silonitis in uisione quo-

9:30 And Solomon reigned in Jerusalem over all Israel que Iaddo uidentis contra Hieroboam filium Nabath 9:30

forty years.

regnauit autem Salomon in Hierusalem super omnem Isra-

9:31 And Solomon slept with his fathers, and he was hel quadraginta annis 9:31 dormiuitque cum patribus suis

buried in the city of David his father: and Rehoboam his son reigned in his stead.

et sepelierunt eum in ciuitate Dauid regnauitque pro eo Roboam filius eius

10:1 And Rehoboam went to Shechem: for to Shechem 10:1 profectus est autem Roboam in Sychem illuc enim

were all Israel come to make him king.

10:2 And it came to pass, when Jeroboam the son of cunctus Israhel conuenerat ut constituerent eum regem 10:2

Nebat, who was in Egypt, whither he fled from the presence of Solomon the king, heard it, that Jeroboam returned

out of Egypt. quod cum audisset Hieroboam filius Nabath qui erat in Ae-

gypto fugerat quippe illuc ante Salomonem statim reuersus

10:3 And they sent and called him. So Jeroboam and all est 10:3 uocaueruntque eum et uenit cum uniuerso Israhel

Israel came and spake to Rehoboam, saying,

10:4 Thy father made our yoke grievous: now therefore et locuti sunt ad Roboam dicentes 10:4 pater tuus durissimo

ease thou somewhat the grievous servitude of thy father, and his heavy yoke that he put upon us, and we will serve

thee. iugo nos pressit tu leuiora impera patre tuo qui nobis grauem

inposuit seruitutem et paululum de onere subleua ut seruia-

10:5 And he said unto them, Come again unto me after mus tibi 10:5 qui ait post tres dies reuertimini ad me cum-

three days. And the people departed.

10:6 And king Rehoboam took counsel with the old men que abisset populus 10:6 iniit consilium cum senibus qui

that had stood before Solomon his father while he yet lived, saying, What counsel give ye me to return answer

to this people? steterant coram patre eius Salomone dum aduiueret dicens 10:7 And they spake unto him, saying, If thou be kind to quid datis consilii ut respondeam populo 10:7 qui dixerunt

this people, and please them, and speak good words to them, they will be thy servants for ever.

ei si placueris populo huic et lenieris eos uerbis clementibus

10:8 But he forsook the counsel which the old men gave seruient tibi omni tempore 10:8 at ille reliquit consilium se-

him, and took counsel with the young men that were brought up with him, that stood before him.

Liber II Paralipomenon

num et cum iuuenibus tractare coepit qui cum eo nutriti fuerant et erant in comitatu illius 10:9 dixitque ad eos quid uo- 10:9 And he said unto them, What advice give ye that we

may return answer to this people, which have spoken to me, saying, Ease somewhat the yoke that thy father did

bis uidetur uel respondere quid debeo populo huic qui dixit put upon us? mihi subleua iugum quod inposuit nobis pater tuus 10:10 10:10 And the young men that were brought up with him

spake unto him, saying, Thus shalt thou answer the people that spake unto thee, saying, Thy father made our yoke heavy, but make thou it somewhat lighter for us; thus shalt thou say unto them, My little finger shall be thicker than my father’s loins.

at illi responderunt ut iuuenes et nutriti cum eo in deliciis atque dixerunt sic loqueris populo qui dixit tibi pater tuus adgrauauit iugum nostrum tu subleua et sic respondebis eis minimus digitus meus grossior est lumbis patris mei 10:11 10:11 For whereas my father put a heavy yoke upon you, I

will put more to your yoke: my father chastised you with whips, but I will chastise you with scorpions.

pater meus inposuit uobis iugum graue et ego maius pondus adponam pater meus cecidit uos flagellis ego uero caedam scorpionibus 10:12 uenit ergo Hieroboam et uniuersus po- 10:12 So Jeroboam and all the people came to Rehoboam

on the third day, as the king bade, saying, Come again to me on the third day.

pulus ad Roboam die tertio sicut praeceperat eis 10:13 res-

10:13 And the king answered them roughly; and king Rehoboam forsook the counsel of the old men,

ponditque rex dura derelicto consilio seniorum 10:14 locu- _{10:14 And answered them after the advice of the young}

men, saying, My father made your yoke heavy, but I will add thereto: my father chastised you with whips, but I will

tusque est iuxta iuuenum uoluntatem pater meus graue uobis _{chastise you with scorpions.} inposuit iugum quod ego grauius faciam pater meus cecidit uos flagellis ego uero caedam scorpionibus 10:15 et non 10:15 So the king hearkened not unto the people: for the

cause was of God, that the LORD might perform his word, which he spake by the hand of Ahijah the Shilonite to Jer-

adquieuit populi precibus erat enim uoluntatis Dei ut con- oboam the son of Nebat. pleretur sermo eius quem locutus fuerat per manum Ahiae Silonitis ad Hieroboam filium Nabath 10:16 populus autem 10:16 And when all Israel saw that the king would not

hearken unto them, the people answered the king, saying, What portion have we in David? and we have none inheritance in the son of Jesse: every man to your tents, O Israel: and now, David, see to thine own house. So all Israel went to their tents.

uniuersus rege duriora dicente sic locutus est ad eum non est nobis pars in Dauid neque hereditas in filio Isai reuertere in tabernacula tua Israhel tu autem pasce domum tuam Dauid et abiit Israhel in tabernacula sua 10:17 super filios 10:17 But as for the children of Israel that dwelt in the

cities of Judah, Rehoboam reigned over them.

autem Israhel qui habitabant in ciuitatibus Iuda regnauit Roboam 10:18 misitque rex Roboam Aduram qui praeerat tri- 10:18 Then king Rehoboam sent Hadoram that was over

the tribute; and the children of Israel stoned him with stones, that he died. But king Rehoboam made speed to

butis et lapidauerunt eum filii Israhel et mortuus est porro get him up to his chariot, to flee to Jerusalem. rex Roboam currum festinauit ascendere et fugit in Hierusa-

lem 10:19 recessitque Israhel a domo Dauid usque ad diem 10:19 And Israel rebelled against the house of David unto

this day.

hanc

11:1 uenit autem Roboam in Hierusalem et conuocauit 11:1 And when Rehoboam was come to Jerusalem, he

gathered of the house of Judah and Benjamin an hundred and fourscore thousand chosen men, which were warriors, to fight against Israel, that he might bring the kingdom again to Rehoboam.

uniuersam domum Iuda et Beniamin in centum octoginta milibus electorum atque bellantium ut dimicaret contra Israhel et conuerteret ad se regnum suum 11:2 factusque est 11:2 But the word of the LORD came to Shemaiah the

man of God, saying,

Biblia Sacra Vulgata

11:3 Speak unto Rehoboam the son of Solomon, king of sermo Domini ad Semeiam hominem Dei dicens 11:3 lo-Judah, and to all Israel in Judah and Benjamin, saying,

quere ad Roboam filium Salomonis regem Iuda et ad uni-

11:4 Thus saith the LORD, Ye shall not go up, nor fight uersum Israhel qui est in Iuda et Beniamin 11:4 haec dicit

against your brethren: return every man to his house: for this thing is done of me. And they obeyed the words of the

LORD, and returned from going against Jeroboam. Dominus non ascendetis neque pugnabitis contra fratres uestros reuertatur unusquisque in domum suam quia mea hoc gestum est uoluntate qui cum audissent sermonem Domini

11:5 And Rehoboam dwelt in Jerusalem, and built cities reuersi sunt nec perrexerunt contra Hieroboam 11:5 habita-

for defence in Judah.

uit autem Roboam in Hierusalem et aedificauit ciuitates mu-

11:6 He built even Bethlehem, and Etam, and Tekoa, ratas in Iuda 11:6 extruxitque Bethleem et Aetham et The-11:7 And Bethzur, and Shoco, and Adullam, cue 11:7 Bethsur quoque et Soccho et Odollam 11:8 nec-

11:8 And Gath, and Mareshah, and Ziph,

11:9 And Adoraim, and Lachish, and Azekah, non Geth et Maresa et Ziph 11:9 sed et Aduram et Lachis et 11:10 And Zorah, and Aijalon, and Hebron, which are in Azecha 11:10 Saraa quoque et Ahilon et Hebron quae erant

Judah and in Benjamin fenced cities.

11:11 And he fortified the strong holds, and put captains in Iuda et Beniamin ciuitates munitissimas 11:11 cumque

in them, and store of victual, and of oil and wine.

clausisset eas muris posuit in eis principes ciborumque hor-

11:12 And in every several city he put shields and spears, rea hoc est olei et uini 11:12 sed et in singulis urbibus fecit

and made them exceeding strong, having Judah and Benjamin on his side.

armamentaria scutorum et hastarum firmauitque eas multa

11:13 And the priests and the Levites that were in all Israel diligentia et imperauit super Iudam et Beniamin 11:13 saresorted to him out of all their coasts.

cerdotes autem et Leuitae qui erant in uniuerso Israhel ue-

11:14 For the Levites left their suburbs and their posses- nerunt ad eum de cunctis sedibus suis 11:14 relinquentes sion, and came to Judah and Jerusalem: for Jeroboam and his sons had cast them off from executing the priest’s of-

fice unto the LORD: suburbana et possessiones suas et transeuntes ad Iudam et

Hierusalem eo quod abiecisset eos Hieroboam et posteri eius

11:15 And he ordained him priests for the high places, and ne sacerdotio Domini fungerentur 11:15 qui constituit sibi

for the devils, and for the calves which he had made.

sacerdotes excelsorum et daemonum uitulorumque quos fe-

11:16 And after them out of all the tribes of Israel such cerat 11:16 sed et de cunctis tribubus Israhel quicumque

as set their hearts to seek the LORD God of Israel came to Jerusalem, to sacrifice unto the LORD God of their fa-

thers. dederant cor suum ut quaererent Dominum Deum Israhel

uenerunt Hierusalem ad immolandas uictimas Domino Deo

11:17 So they strengthened the kingdom of Judah, and patrum suorum 11:17 et roborauerunt regnum Iuda et con-

made Rehoboam the son of Solomon strong, three years: for three years they walked in the way of David and

Solomon. firmauerunt Roboam filium Salomonis per tres annos am-

bulauerunt enim in uiis Dauid et Salomonis annis tantum

11:18 And Rehoboam took him Mahalath the daughter of tribus 11:18 duxit autem Roboam uxorem Maalath filiam

Jerimoth the son of David to wife, and Abihail the daughter of Eliab the son of Jesse;

Hierimuth filii Dauid Abiail quoque filiam Heliab filii Isai

11:19 Which bare him children; Jeush, and Shamariah, 11:19 quae peperit ei filios Ieus et Somoriam et Zoom 11:20

and Zaham. 11:20 And after her he took Maachah the daughter of Absalom; which bare him Abijah, and Attai, and Ziza, and Shelomith.

post hanc quoque accepit Maacha filiam Absalom quae pe-

Liber II Paralipomenon

perit ei Abia et Ethai et Ziza et Salumith 11:21 amauit au- 11:21 And Rehoboam loved Maachah the daughter of Ab-

salom above all his wives and his concubines: (for he took eighteen wives, and threescore concubines; and be-

tem Roboam Maacha filiam Absalom super omnes uxores gat twenty and eight sons, and threescore daughters.)

suas et concubinas nam uxores decem et octo duxerat concubinasque sexaginta et genuit uiginti octo filios et sexaginta filias 11:22 constituit uero in capite Abiam filium Maacha 11:22 And Rehoboam made Abijah the son of Maachah

the chief, to be ruler among his brethren: for he thought to make him king.

ducem super fratres suos ipsum enim regem facere cogitabat 11:23 qui sapientior fuit et potentior super omnes filios eius 11:23 And he dealt wisely, and dispersed of all his children

throughout all the countries of Judah and Benjamin, unto every fenced city: and he gave them victual in abundance.

et in cunctis finibus Iuda et Beniamin et in uniuersis ciuitati- And he desired many wives. bus muratis praebuitque eis escas plurimas et multas petiuit uxores

12:1 cumque roboratum fuisset regnum Roboam et con- 12:1 And it came to pass, when Rehoboam had established the kingdom, and had strengthened himself, he forsook the law of the LORD, and all Israel with him.

fortatum dereliquit legem Domini et omnis Israhel cum eo 12:2 anno autem quinto regni Roboam ascendit Sesac rex 12:2 And it came to pass, that in the fifth year of king Re-

hoboam Shishak king of Egypt came up against Jerusalem, because they had transgressed against the LORD,

Aegypti in Hierusalem quia peccauerunt Domino 12:3 cum

12:3 With twelve hundred chariots, and threescore thousand horsemen: and the people were without number that came with him out of Egypt; the Lubims, the Sukkiims, and the Ethiopians.

mille ducentis curribus et sexaginta milibus equitum nec erat numerus uulgi quod uenerat cum eo ex Aegypto Lybies scilicet et Trogoditae et Aethiopes 12:4 cepitque ciuitates mu- 12:4 And he took the fenced cities which pertained to Ju-

dah, and came to Jerusalem.

nitissimas in Iuda et uenit usque Hierusalem 12:5 Semeias 12:5 Then came Shemaiah the prophet to Rehoboam, and

to the princes of Judah, that were gathered together to Jerusalem because of Shishak, and said unto them, Thus saith the LORD, Ye have forsaken me, and therefore have I also left you in the hand of Shishak.

autem propheta ingressus est ad Roboam et principes Iuda qui congregati fuerant in Hierusalem fugientes Sesac dixitque ad eos haec dicit Dominus uos reliquistis me et ego reliqui uos in manu Sesac 12:6 consternatique principes Isra- 12:6 Whereupon the princes of Israel and the king hum-

bled themselves; and they said, The LORD is righteous.

hel et rex dixerunt iustus est Dominus 12:7 cumque uidisset 12:7 And when the LORD saw that they humbled them-

selves, the word of the LORD came to Shemaiah, saying, They have humbled themselves; therefore I will not destroy them, but I will grant them some deliverance; and my wrath shall not be poured out upon Jerusalem by the hand of Shishak.

Dominus quod humiliati essent factus est sermo Domini ad Semeiam dicens quia humiliati sunt non disperdam eos daboque eis pauxillum auxilii et non stillabit furor meus super

Hierusalem per manum Sesac 12:8 uerumtamen seruient ei 12:8 Nevertheless they shall be his servants; that they may know my service, and the service of the kingdoms of the countries.

ut sciant distantiam seruitutis meae et seruitutis regni terrarum 12:9 recessit itaque Sesac rex Aegypti ab Hierusalem 12:9 So Shishak king of Egypt came up against Jerusalem,

and took away the treasures of the house of the LORD, and the treasures of the king’s house; he took all: he carried

sublatis thesauris domus Domini et domus regis omniaque away also the shields of gold which Solomon had made. secum tulit et clypeos aureos quos fecerat Salomon 12:10 12:10 Instead of which king Rehoboam made shields of

brass, and committed them to the hands of the chief of the guard, that kept the entrance of the king’s house.

pro quibus fecit rex aeneos et tradidit illos principibus scutariorum qui custodiebant uestibulum palatii 12:11 cumque 12:11 And when the king entered into the house of the

LORD, the guard came and fetched them, and brought them again into the guard chamber.

Biblia Sacra Vulgata

introiret rex domum Domini ueniebant scutarii et tollebant

12:12 And when he humbled himself, the wrath of the eos iterumque referebant ad armamentarium suum 12:12

LORD turned from him, that he would not destroy him altogether: and also in Judah things went well.

uerumtamen quia humiliati sunt auersa est ab eis ira Domini nec deleti sunt penitus siquidem et in Iuda inuenta sunt

12:13 So king Rehoboam strengthened himself in opera bona 12:13 confortatus est igitur rex Roboam in Hie-

Jerusalem, and reigned: for Rehoboam was one and forty years old when he began to reign, and he reigned seventeen years in Jerusalem, the city which the LORD had chosen out of all the tribes of Israel, to put his name there. And his mother’s name was Naamah an Ammonitess.

rusalem atque regnauit quadraginta autem et unius anni erat cum regnare coepisset et decem septemque annis regnauit in Hierusalem urbe quam elegit Dominus ut confirmaret nomen suum ibi de cunctis tribubus Israhel nomenque matris eius

12:14 And he did evil, because he prepared not his heart Naama Ammanitis 12:14 fecit autem malum et non praeto seek the LORD.

12:15 Now the acts of Rehoboam, first and last, are they parauit cor suum ut quaereret Dominum 12:15 opera uero

not written in the book of Shemaiah the prophet, and of Iddo the seer concerning genealogies? And there were

wars between Rehoboam and Jeroboam continually. Roboam prima et nouissima scripta sunt in libris Semeiae prophetae et Addo uidentis et diligenter exposita pugnaueruntque aduersum se Roboam et Hieroboam cunctis diebus

12:16 And Rehoboam slept with his fathers, and was 12:16 et dormiuit Roboam cum patribus suis sepultusque est

buried in the city of David: and Abijah his son reigned in his stead.

in ciuitate Dauid et regnauit Abia filius eius pro eo

13:1 Now in the eighteenth year of king Jeroboam began 13:1 anno octauodecimo regis Hieroboam regnauit Abia

Abijah to reign over Judah.

13:2 He reigned three years in Jerusalem. His mother’s super Iudam 13:2 tribus annis regnauit in Hierusalem no-

name also was Michaiah the daughter of Uriel of Gibeah. And there was war between Abijah and Jeroboam.

menque matris eius Michaia filia Vrihel de Gabaa et erat bel-

13:3 And Abijah set the battle in array with an army of lum inter Abia et Hieroboam 13:3 cumque inisset Abia cer-

valiant men of war, even four hundred thousand chosen men: Jeroboam also set the battle in array against him with eight hundred thousand chosen men, being mighty men of valour.

tamen et haberet bellicosissimos uiros et electorum quadringenta milia Hieroboam instruxit e contra aciem octingenta milia uirorum qui et ipsi electi erant et ad bella fortissimi

13:4 And Abijah stood up upon mount Zemaraim, which 13:4 stetit igitur Abia super montem Someron qui erat in

is in mount Ephraim, and said, Hear me, thou Jeroboam, and all Israel;

Ephraim et ait audi Hieroboam et omnis Israhel 13:5 num

13:5 Ought ye not to know that the LORD God of Israel gave the kingdom over Israel to David for ever, even to him and to his sons by a covenant of salt?

ignoratis quod Dominus Deus Israhel dederit regnum Dauid super Israhel in sempiternum ipsi et filiis eius pactum salis

13:6 Yet Jeroboam the son of Nebat, the servant of 13:6 et surrexit Hieroboam filius Nabath seruus Salomonis

Solomon the son of David, is risen up, and hath rebelled against his lord.

filii Dauid et rebellauit contra dominum suum 13:7 congre-

13:7 And there are gathered unto him vain men, the children of Belial, and have strengthened themselves against Rehoboam the son of Solomon, when Rehoboam was young and tenderhearted, and could not withstand them.

gatique sunt ad eum uiri uanissimi et filii Belial et praeualuerunt contra Roboam filium Salomonis porro Roboam erat

13:8 And now ye think to withstand the kingdom of the rudis et corde pauido nec potuit resistere eis 13:8 nunc ergo

LORD in the hand of the sons of David; and ye be a great multitude, and there are with your golden calves, which

Jeroboam made you for gods. uos dicitis quod resistere possitis regno Domini quod pos-

Liber II Paralipomenon

sidet per filios Dauid habetisque grandem populi multitudinem atque uitulos aureos quos fecit uobis Hieroboam in deos 13:9 et eiecistis sacerdotes Domini filios Aaron atque Leui- 13:9 Have ye not cast out the priests of the LORD, the sons

of Aaron, and the Levites, and have made you priests after the manner of the nations of other lands? so that whosoever cometh to consecrate himself with a young bullock and seven rams, the same may be a priest of them that are no gods.

tas et fecistis uobis sacerdotes sicut omnes populi terrarum quicumque uenerit et initiauerit manum suam in tauro in bubus et in arietibus septem fit sacerdos eorum qui non sunt dii 13:10 noster autem Dominus Deus est quem non relinquimus 13:10 But as for us, the LORD is our God, and we have

not forsaken him; and the priests, which minister unto the LORD, are the sons of Aaron, and the Levites wait upon

sacerdotesque ministrant Domino de filiis Aaron et Leuitae their business: sunt in ordine suo 13:11 holocausta quoque offerunt Do- 13:11 And they burn unto the LORD every morning and

every evening burnt sacrifices and sweet incense: the shewbread also set they in order upon the pure table; and the candlestick of gold with the lamps thereof, to burn every evening: for we keep the charge of the LORD our God; but ye have forsaken him.

mino per singulos dies mane et uespere et thymiama iuxta legis praecepta confectum et proponuntur panes in mensa mundissima estque apud nos candelabrum aureum et lucernae eius ut accendantur semper ad uesperam nos quippe custodimus praecepta Domini Dei nostri quem uos reliquistis 13:12 ergo in exercitu nostro dux Deus est et sacerdotes eius 13:12 And, behold, God himself is with us for our captain,

and his priests with sounding trumpets to cry alarm against you. O children of Israel, fight ye not against the LORD

qui clangunt tubis et resonant contra uos filii Israhel nolite God of your fathers; for ye shall not prosper. pugnare contra Dominum Deum patrum uestrorum quia non uobis expedit 13:13 haec illo loquente Hieroboam retro mo- 13:13 But Jeroboam caused an ambushment to come about

behind them: so they were before Judah, and the ambushment was behind them.

liebatur insidias cumque ex aduerso hostium staret ignorantem Iudam suo ambiebat exercitu 13:14 respiciensque Iu- 13:14 And when Judah looked back, behold, the battle was

before and behind: and they cried unto the LORD, and the priests sounded with the trumpets.

das uidit instare bellum ex aduerso et post tergum et clamauit ad Dominum ac sacerdotes tubis canere coeperunt 13:15 13:15 Then the men of Judah gave a shout: and as the

men of Judah shouted, it came to pass, that God smote Jeroboam and all Israel before Abijah and Judah.

omnesque uiri Iuda uociferati sunt et ecce illis clamantibus perterruit Deus Hieroboam et omnem Israhel qui stabat ex aduerso Abia et Iuda 13:16 fugeruntque filii Israhel Iudam 13:16 And the children of Israel fled before Judah: and

God delivered them into their hand.

et tradidit eos Deus in manu eorum 13:17 percussit ergo eos 13:17 And Abijah and his people slew them with a great

slaughter: so there fell down slain of Israel five hundred thousand chosen men.

Abia et populus eius plaga magna et corruerunt uulnerati ex Israhel quingenta milia uirorum fortium 13:18 humiliatique 13:18 Thus the children of Israel were brought under at

that time, and the children of Judah prevailed, because they relied upon the LORD God of their fathers.

sunt filii Israhel in tempore illo et uehementissime confortati filii Iuda eo quod sperassent in Domino Deo patrum suorum 13:19 persecutus est autem Abia fugientem Hieroboam et ce- 13:19 And Abijah pursued after Jeroboam, and took cities

from him, Bethel with the towns thereof, and Jeshanah with the towns thereof, and Ephraim with the towns

pit ciuitates eius Bethel et filias eius et Hiesena cum filiabus thereof. suis Ephron quoque et filias eius 13:20 nec ualuit ultra re- 13:20 Neither did Jeroboam recover strength again in the

days of Abijah: and the LORD struck him, and he died.

Biblia Sacra Vulgata

sistere Hieroboam in diebus Abia quem percussit Dominus

13:21 But Abijah waxed mighty, and married fourteen et mortuus est 13:21 igitur Abia confortato imperio suo ac-

wives, and begat twenty and two sons, and sixteen daughters.

cepit uxores quattuordecim procreauitque uiginti duos filios

13:22 And the rest of the acts of Abijah, and his ways, and et sedecim filias 13:22 reliqua autem sermonum Abia uiahis sayings, are written in the story of the prophet Iddo.

rumque et operum eius scripta sunt diligentissime in libro prophetae Addo

14:1 So Abijah slept with his fathers, and they buried him 14:1 dormiuit autem Abia cum patribus suis et sepelierunt

in the city of David: and Asa his son reigned in his stead. In his days the land was quiet ten years.

eum in ciuitate Dauid regnauitque Asa filius eius pro eo in

14:2 And Asa did that which was good and right in the cuius diebus quieuit terra annis decem 14:2 fecit autem Asa

eyes of the LORD his God:

quod bonum et placitum erat in conspectu Dei sui et subuer-

14:3 For he took away the altars of the strange gods, and tit altaria peregrini cultus et excelsa 14:3 et confregit sta-

the high places, and brake down the images, and cut down the groves:

tuas lucosque succidit 14:4 ac praecepit Iudae ut quaereret

14:4 And commanded Judah to seek the LORD God of their fathers, and to do the law and the commandment.

Dominum Deum patrum suorum et faceret legem et uniuersa

14:5 Also he took away out of all the cities of Judah the mandata 14:5 et abstulit e cunctis urbibus Iuda aras et fana

high places and the images: and the kingdom was quiet before him.

et regnauit in pace 14:6 aedificauit quoque urbes munitas

14:6 And he built fenced cities in Judah: for the land had rest, and he had no war in those years; because the LORD had given him rest.

in Iuda quia quietus erat et nulla temporibus eius bella sur-

14:7 Therefore he said unto Judah, Let us build these rexerant pacem Domino largiente 14:7 dixit autem Iudae

cities, and make about them walls, and towers, gates, and bars, while the land is yet before us; because we have sought the LORD our God, we have sought him, and he hath given us rest on every side. So they built and prospered.

aedificemus ciuitates istas et uallemus muris et roboremus turribus et portis et seris donec a bellis quieta sunt omnia eo quod quaesierimus Dominum Deum patrum nostrorum et dederit nobis pacem per gyrum aedificauerunt igitur et nul-

14:8 And Asa had an army of men that bare targets and lum in extruendo inpedimentum fuit 14:8 habuit autem Asa

spears, out of Judah three hundred thousand; and out of Benjamin, that bare shields and drew bows, two hundred and fourscore thousand: all these were mighty men of valour.

in exercitu suo portantium scuta et hastas de Iuda trecenta milia de Beniamin uero scutariorum et sagittariorum ducenta

14:9 And there came out against them Zerah the Ethiopian octoginta milia omnes isti uiri fortissimi 14:9 egressus est

with an host of a thousand thousand, and three hundred chariots; and came unto Mareshah.

autem contra eos Zara Aethiops cum exercitu decies cen-

14:10 Then Asa went out against him, and they set the tena milia et curribus trecentis et uenit usque Maresa 14:10

battle in array in the valley of Zephathah at Mareshah.

porro Asa perrexit obuiam et instruxit aciem ad bellum in

14:11 And Asa cried unto the LORD his God, and said, ualle Sephata quae est iuxta Maresa 14:11 et inuocauit Do-

LORD, it is nothing with thee to help, whether with many, or with them that have no power: help us, O LORD our God; for we rest on thee, and in thy name we go against this multitude. O LORD, thou art our God; let no man prevail against thee.

minum Deum et ait Domine non est apud te ulla distantia utrum in paucis auxilieris an in pluribus adiuua nos Domine Deus noster in te enim et in tuo nomine habentes fiduciam uenimus contra hanc multitudinem Domine Deus noster tu

Liber II Paralipomenon

es non praeualeat contra te homo 14:12 exterruit itaque Do- 14:12 So the LORD smote the Ethiopians before Asa, and

before Judah; and the Ethiopians fled.

minus Aethiopas coram Asa et Iuda fugeruntque Aethiopes 14:13 et persecutus est eos Asa et populus qui cum eo erat us- 14:13 And Asa and the people that were with him pursued

them unto Gerar: and the Ethiopians were overthrown, that they could not recover themselves; for they were destroyed before the LORD, and before his host; and they carried away very much spoil.

que Gerar et ruerunt Aethiopes usque ad internicionem quia Domino caedente contriti sunt et exercitu illius proeliante tulerunt ergo spolia multa 14:14 et percusserunt omnes ci- 14:14 And they smote all the cities round about Gerar; for

the fear of the LORD came upon them: and they spoiled all the cities; for there was exceeding much spoil in them.

uitates per circuitum Gerare grandis quippe cunctos terror inuaserat et diripuerunt urbes et multam praedam asportauerunt 14:15 sed et caulas ouium destruentes tulerunt pe- 14:15 They smote also the tents of cattle, and carried away

sheep and camels in abundance, and returned to Jerusalem.

corum infinitam multitudinem et camelorum reuersique sunt Hierusalem

15:1 Azarias autem filius Oded facto in se spiritu Dei 15:1 And the Spirit of God came upon Azariah the son of

Oded:

15:2 egressus est in occursum Asa et dixit ei audite me Asa 15:2 And he went out to meet Asa, and said unto him,

Hear ye me, Asa, and all Judah and Benjamin; The LORD is with you, while ye be with him; and if ye seek him, he will be found of you; but if ye forsake him, he will forsake you.

et omnis Iuda et Beniamin Dominus uobiscum quia fuistis cum eo si quaesieritis eum inuenietis si autem dereliqueritis derelinquet uos 15:3 transibunt autem multi dies in Israhel 15:3 Now for a long season Israel hath been without the

true God, and without a teaching priest, and without law.

absque Deo uero et absque sacerdote doctore et absque lege 15:4 cumque reuersi fuerint in angustia sua ad Dominum De- 15:4 But when they in their trouble did turn unto the

LORD God of Israel, and sought him, he was found of them.

um Israhel et quaesierint eum repperient 15:5 in tempore

15:5 And in those times there was no peace to him that went out, nor to him that came in, but great vexations were upon all the inhabitants of the countries.

illo non erit pax egredienti et ingredienti sed terrores undique in cunctis habitatoribus terrarum 15:6 pugnabit enim 15:6 And nation was destroyed of nation, and city of city:

for God did vex them with all adversity.

gens contra gentem et ciuitas contra ciuitatem quia Dominus conturbabit eos in omni angustia 15:7 uos ergo confor- 15:7 Be ye strong therefore, and let not your hands be

weak: for your work shall be rewarded.

tamini et non dissoluantur manus uestrae erit enim merces operi uestro 15:8 quod cum audisset Asa uerba scilicet et 15:8 And when Asa heard these words, and the prophecy

of Oded the prophet, he took courage, and put away the abominable idols out of all the land of Judah and Benjamin, and out of the cities which he had taken from mount Ephraim, and renewed the altar of the LORD, that was before the porch of the LORD.

prophetiam Oded prophetae confortatus est et abstulit idola de omni terra Iuda et Beniamin et ex urbibus quas ceperat montis Ephraim et dedicauit altare Domini quod erat ante porticum Domini 15:9 congregauitque uniuersum Iuda et 15:9 And he gathered all Judah and Benjamin, and the

strangers with them out of Ephraim and Manasseh, and out of Simeon: for they fell to him out of Israel in abundance, when they saw that the LORD his God was with him.

Beniamin et aduenas cum eis de Ephraim et de Manasse et de Symeon plures enim ad eum confugerant ex Israhel ui-

dentes quod Dominus Deus illius esset cum eo 15:10 cum- 15:10 So they gathered themselves together at Jerusalem

in the third month, in the fifteenth year of the reign of Asa.

que uenissent Hierusalem mense tertio anno quintodecimo

Biblia Sacra Vulgata

15:11 And they offered unto the LORD the same time, of regni Asa 15:11 immolauerunt Domino in die illa de manu-

the spoil which they had brought, seven hundred oxen and seven thousand sheep.

biis et praeda quam adduxerant boues septingentos et arie-

15:12 And they entered into a covenant to seek the LORD tes septem milia 15:12 et intrauit ex more ad corroboran-

God of their fathers with all their heart and with all their soul;

dum foedus ut quaererent Dominum Deum patrum suorum

15:13 That whosoever would not seek the LORD God in toto corde et in tota anima sua 15:13 si quis autem inquit

of Israel should be put to death, whether small or great, whether man or woman.

non quaesierit Dominum Deum Israhel moriatur a minimo

15:14 And they sware unto the LORD with a loud voice, usque ad maximum a uiro usque ad mulierem 15:14 iuraue-

and with shouting, and with trumpets, and with cornets.

runtque Domino uoce magna in iubilo et in clangore tubae

15:15 And all Judah rejoiced at the oath: for they had et in sonitu bucinarum 15:15 omnes qui erant in Iuda cum

sworn with all their heart, and sought him with their whole desire; and he was found of them: and the LORD gave

them rest round about. execratione in omni enim corde suo iurauerunt et in tota uo-

luntate quaesierunt eum et inuenerunt praestititque eis Do-

15:16 And also concerning Maachah the mother of Asa minus requiem per circuitum 15:16 sed et Maacham mat-

the king, he removed her from being queen, because she had made an idol in a grove: and Asa cut down her idol,

and stamped it, and burnt it at the brook Kidron. rem Asa regis ex augusto deposuit imperio eo quod fecisset

in luco simulacrum Priapi quod omne contriuit et in frusta

15:17 But the high places were not taken away out of Is- comminuens conbusit in torrente Cedron 15:17 excelsa au-

rael: nevertheless the heart of Asa was perfect all his days.

tem derelicta sunt in Israhel attamen cor Asa erat perfectum

15:18 And he brought into the house of God the things cunctis diebus eius 15:18 ea quae uouerat pater suus et ipse

that his father had dedicated, and that he himself had dedicated, silver, and gold, and vessels.

intulit in domum Domini argentum et aurum uasorumque di-

15:19 And there was no more war unto the five and thirti- uersam supellectilem 15:19 bellum uero non fuit usque ad

eth year of the reign of Asa.

tricesimum quintum annum regni Asa

16:1 In the six and thirtieth year of the reign of Asa Baasha 16:1 anno autem tricesimo sexto regni eius ascendit Baking of Israel came up against Judah, and built Ramah, to the intent that he might let none go out or come in to Asa

king of Judah. asa rex Israhel in Iudam et muro circumdabat Rama ut nul-16:2 Then Asa brought out silver and gold out of the trea- lus tute posset egredi et ingredi de regno Asa 16:2 protulit

sures of the house of the LORD and of the king’s house, and sent to Benhadad king of Syria, that dwelt at Damas-

cus, saying, ergo Asa argentum et aurum de thesauris domus Domini et

de thesauris regis misitque ad Benadad regem Syriae qui ha-

16:3 There is a league between me and thee, as there was bitabat in Damasco dicens 16:3 foedus inter me et te est pa-

between my father and thy father: behold, I have sent thee silver and gold; go, break thy league with Baasha king of

Israel, that he may depart from me. ter quoque meus et pater tuus habuere concordiam quam ob

rem misi tibi argentum et aurum ut rupto foedere quod habes

16:4 And Benhadad hearkened unto king Asa, and sent cum Baasa rege Israhel facias eum a me recedere 16:4 quo

the captains of his armies against the cities of Israel; and they smote Ijon, and Dan, and Abelmaim, and all the store

cities of Naphtali. conperto Benadad misit principes exercituum suorum ad ur-

bes Israhel qui percusserunt Ahion et Dan et Abelmaim et

16:5 And it came to pass, when Baasha heard it, that he uniuersas urbes muratas Nepthalim 16:5 quod cum audis-

left off building of Ramah, and let his work cease.

set Baasa desiuit aedificare Rama et intermisit opus suum

Liber II Paralipomenon

16:6 porro Asa rex adsumpsit uniuersum Iudam et tulerunt 16:6 Then Asa the king took all Judah; and they carried

away the stones of Ramah, and the timber thereof, wherewith Baasha was building; and he built therewith Geba and

lapides Rama et ligna quae aedificationi praeparauerat Ba- Mizpah. asa aedificauitque ex eis Gabaa et Maspha 16:7 in tempore 16:7 And at that time Hanani the seer came to Asa king

of Judah, and said unto him, Because thou hast relied on the king of Syria, and not relied on the LORD thy God, therefore is the host of the king of Syria escaped out of thine hand.

illo uenit Anani propheta ad Asam regem Iuda et dixit ei quia habuisti fiduciam in rege Syriae et non in Domino Deo tuo idcirco euasit Syriae regis exercitus de manu tua 16:8 16:8 Were not the Ethiopians and the Lubims a huge host,

with very many chariots and horsemen? yet, because thou didst rely on the LORD, he delivered them into thine hand.

nonne Aethiopes et Lybies multo plures erant quadrigis et equitibus et multitudine nimia quos cum Domino credidisses tradidit in manu tua 16:9 oculi enim eius contemplan- 16:9 For the eyes of the LORD run to and fro throughout

the whole earth, to shew himself strong in the behalf of them whose heart is perfect toward him. Herein thou hast done foolishly: therefore from henceforth thou shalt have wars.

tur uniuersam terram et praebent fortitudinem his qui corde perfecto credunt in eum stulte igitur egisti et propter hoc ex praesenti tempore contra te bella consurgent 16:10 iratus- 16:10 Then Asa was wroth with the seer, and put him in a

prison house; for he was in a rage with him because of this thing. And Asa oppressed some of the people the same

que Asa aduersus uidentem iussit eum mitti in neruum ualde time. quippe super hoc fuerat indignatus et interfecit de populo

in tempore illo plurimos 16:11 opera autem Asa prima et 16:11 And, behold, the acts of Asa, first and last, lo, they

are written in the book of the kings of Judah and Israel.

nouissima scripta sunt in libro regum Iuda et Israhel 16:12 16:12 And Asa in the thirty and ninth year of his reign was

diseased in his feet, until his disease was exceeding great: yet in his disease he sought not to the LORD, but to the

aegrotauit etiam Asa anno tricesimo nono regni sui dolore physicians. pedum uehementissimo et nec in infirmitate sua quaesiuit

Dominum sed magis in medicorum arte confisus est 16:13 16:13 And Asa slept with his fathers, and died in the one

and fortieth year of his reign.

dormiuitque cum patribus suis et mortuus est anno quadragesimo primo regni sui 16:14 et sepelierunt eum in sepulchro 16:14 And they buried him in his own sepulchres, which

he had made for himself in the city of David, and laid him in the bed which was filled with sweet odours and divers kinds of spices prepared by the apothecaries’ art: and they made a very great burning for him.

suo quod foderat sibi in ciuitate Dauid posueruntque eum super lectulum suum plenum aromatibus et unguentis meretriciis quae erant pigmentariorum arte confecta et conbuserunt super eum ambitione nimia

17:1 regnauit autem Iosaphat filius eius pro eo et inualuit 17:1 And Jehoshaphat his son reigned in his stead, and

strengthened himself against Israel.

contra Israhel 17:2 constituitque militum numeros in cunc- 17:2 And he placed forces in all the fenced cities of Judah,

and set garrisons in the land of Judah, and in the cities of Ephraim, which Asa his father had taken.

tis urbibus Iudae quae erant uallatae muris praesidiaque disposuit in terra Iuda et in ciuitatibus Ephraim quas ceperat Asa pater eius 17:3 et fuit Dominus cum Iosaphat quia am- 17:3 And the LORD was with Jehoshaphat, because he

walked in the first ways of his father David, and sought not unto Baalim;

bulauit in uiis Dauid patris sui primis et non sperauit in Baalim 17:4 sed in Deo patris sui et perrexit in praeceptis illius 17:4 But sought to the Lord God of his father, and walked

in his commandments, and not after the doings of Israel.

et non iuxta peccata Israhel 17:5 confirmauitque Dominus 17:5 Therefore the LORD stablished the kingdom in his

hand; and all Judah brought to Jehoshaphat presents; and he had riches and honour in abundance.

Biblia Sacra Vulgata

regnum in manu eius et dedit omnis Iuda munera Iosaphat

17:6 And his heart was lifted up in the ways of the LORD: factaeque sunt ei infinitae diuitiae et multa gloria 17:6 cum-

moreover he took away the high places and groves out of Judah.

que sumpsisset cor eius audaciam propter uias Domini etiam

17:7 Also in the third year of his reign he sent to excelsa et lucos de Iuda abstulit 17:7 tertio autem anno re-

his princes, even to Benhail, and to Obadiah, and to Zechariah, and to Nethaneel, and to Michaiah, to teach

in the cities of Judah. gni sui misit de principibus suis Benail et Obdiam et Zac-

chariam et Nathanahel et Micheam ut docerent in ciuitati-

17:8 And with them he sent Levites, even Shemaiah, bus Iuda 17:8 et cum eis Leuitas Semeiam et Nathaniam

and Nethaniah, and Zebadiah, and Asahel, and Shemiramoth, and Jehonathan, and Adonijah, and Tobijah, and Tobadonijah, Levites; and with them Elishama and Jehoram, priests.

et Zabadiam Asahel quoque et Semiramoth et Ionathan Adoniam et Tobiam et Tobadoniam Leuitas et cum eis Elisama

17:9 And they taught in Judah, and had the book of the et Ioram sacerdotes 17:9 docebantque in Iuda habentes li-

law of the LORD with them, and went about throughout all the cities of Judah, and taught the people.

brum legis Domini et circuibant cunctas urbes Iuda atque

17:10 And the fear of the LORD fell upon all the king- erudiebant populum 17:10 itaque factus est pauor Domini

doms of the lands that were round about Judah, so that they made no war against Jehoshaphat.

super omnia regna terrarum quae erant per gyrum Iuda nec

17:11 Also some of the Philistines brought Jehoshaphat audebant bellare contra Iosaphat 17:11 sed et Philisthei Io-

presents, and tribute silver; and the Arabians brought him flocks, seven thousand and seven hundred rams, and seven

thousand and seven hundred he goats. saphat munera deferebant et uectigal argenti Arabes quoque

adducebant pecora arietum septem milia septingentos et hir-

17:12 And Jehoshaphat waxed great exceedingly; and he cos totidem 17:12 creuit ergo Iosaphat et magnificatus est

built in Judah castles, and cities of store.

usque in sublime atque aedificauit in Iuda domos ad instar

17:13 And he had much business in the cities of Ju- turrium urbesque muratas 17:13 et multa opera patrauit in

dah: and the men of war, mighty men of valour, were in Jerusalem.

urbibus Iuda uiri quoque bellatores et robusti erant in Hie-

17:14 And these are the numbers of them according to the rusalem 17:14 quorum iste numerus per domos atque fami-

house of their fathers: Of Judah, the captains of thousands; Adnah the chief, and with him mighty men of valour three

hundred thousand. lias singulorum in Iuda principes exercitus Ednas dux et cum 17:15 And next to him was Jehohanan the captain, and eo robustissimorum trecenta milia 17:15 post hunc Iohanan

with him two hundred and fourscore thousand.

17:16 And next him was Amasiah the son of Zichri, who princeps et cum eo ducenta octoginta milia 17:16 post is-

willingly offered himself unto the LORD; and with him two hundred thousand mighty men of valour.

tum quoque Amasias filius Zechri consecratus Domino et

17:17 And of Benjamin; Eliada a mighty man of valour, cum eo ducenta milia uirorum fortium 17:17 hunc sequeba-

and with him armed men with bow and shield two hundred thousand.

tur robustus ad proelia Heliada et cum eo tenentium arcum et

17:18 And next him was Jehozabad, and with him an hun- clypeum ducenta milia 17:18 post istum etiam Iozabath et

dred and fourscore thousand ready prepared for the war.

17:19 These waited on the king, beside those whom the cum eo centum octoginta milia expeditorum militum 17:19

king put in the fenced cities throughout all Judah.

hii omnes erant ad manum regis exceptis aliis quos posuerat in urbibus muratis et in uniuerso Iuda

18:1 Now Jehoshaphat had riches and honour in abun- 18:1 fuit ergo Iosaphat diues et inclitus multum et adfini-

dance, and joined affinity with Ahab.

18:2 And after certain years he went down to Ahab to tate coniunctus est Ahab 18:2 descenditque post annos ad

Samaria. And Ahab killed sheep and oxen for him in abundance, and for the people that he had with him, and persuaded him to go up with him to Ramothgilead.

Liber II Paralipomenon

eum in Samariam ad cuius aduentum mactauit Ahab arietes et boues plurimos et populo qui uenerat cum eo persuasitque illi ut ascenderet in Ramoth Galaad 18:3 dixitque Ahab rex 18:3 And Ahab king of Israel said unto Jehoshaphat king

of Judah, Wilt thou go with me to Ramothgilead? And he answered him, I am as thou art, and my people as thy

Israhel ad Iosaphat regem Iuda ueni mecum in Ramoth Ga- people; and we will be with thee in the war. laad cui ille respondit ut ego et tu sicut populus tuus sic et populus meus tecumque erimus in bello 18:4 dixitque Iosa- 18:4 And Jehoshaphat said unto the king of Israel, En-

quire, I pray thee, at the word of the LORD to day.

phat ad regem Israhel consule obsecro inpraesentiarum sermonem Domini 18:5 congregauitque rex Israhel propheta- 18:5 Therefore the king of Israel gathered together of

prophets four hundred men, and said unto them, Shall we go to Ramothgilead to battle, or shall I forbear? And they

rum quadringentos uiros et dixit ad eos in Ramoth Galaad said, Go up; for God will deliver it into the king’s hand. ad bellandum ire debemus an quiescere at illi ascende inquiunt et tradet Deus in manu regis 18:6 dixitque Iosaphat 18:6 But Jehoshaphat said, Is there not here a prophet of

the LORD besides, that we might enquire of him?

numquid non est hic prophetes Domini ut ab illo etiam requiramus 18:7 et ait rex Israhel ad Iosaphat est uir unus a 18:7 And the king of Israel said unto Jehoshaphat, There

is yet one man, by whom we may enquire of the LORD: but I hate him; for he never prophesied good unto me, but always evil: the same is Micaiah the son of Imla. And Jehoshaphat said, Let not the king say so.

quo possumus quaerere Domini uoluntatem sed ego odi eum quia non prophetat mihi bonum sed malum omni tempore est autem Micheas filius Iembla dixitque Iosaphat ne loqua-

ris rex hoc modo 18:8 uocauit ergo rex Israhel unum de 18:8 And the king of Israel called for one of his officers,

and said, Fetch quickly Micaiah the son of Imla.

eunuchis et dixit ei uoca cito Micheam filium Iembla 18:9 18:9 And the king of Israel and Jehoshaphat king of Judah

sat either of them on his throne, clothed in their robes, and they sat in a void place at the entering in of the gate of

porro rex Israhel et Iosaphat rex Iuda uterque sedebant in so- Samaria; and all the prophets prophesied before them. lio suo uestiti cultu regio sedebant autem in area iuxta portam Samariae omnesque prophetae uaticinabantur coram eis 18:10 Sedecias uero filius Chanana fecit sibi cornua ferrea et 18:10 And Zedekiah the son of Chenaanah had made him

horns of iron, and said, Thus saith the LORD, With these thou shalt push Syria until they be consumed.

ait haec dicit Dominus his uentilabis Syriam donec conteras eam 18:11 omnesque prophetae similiter prophetabant 18:11 And all the prophets prophesied so, saying, Go up

to Ramothgilead, and prosper: for the LORD shall deliver it into the hand of the king.

atque dicebant ascende in Ramoth Galaad et prosperaberis et tradet eos Dominus in manu regis 18:12 nuntius autem 18:12 And the messenger that went to call Micaiah spake

to him, saying, Behold, the words of the prophets declare good to the king with one assent; let thy word therefore, I

qui ierat ad uocandum Micheam ait illi en uerba omnium pray thee, be like one of their’s, and speak thou good. prophetarum uno ore bona regi adnuntiant quaeso ergo te ut et sermo tuus ab eis non dissentiat loquarisque prospera 18:13 cui respondit Micheas uiuit Dominus quia quodcum- 18:13 And Micaiah said, As the LORD liveth, even what

my God saith, that will I speak.

que dixerit Deus meus hoc loquar 18:14 uenit ergo ad re- 18:14 And when he was come to the king, the king said

unto him, Micaiah, shall we go to Ramothgilead to battle, or shall I forbear? And he said, Go ye up, and prosper, and

gem cui rex ait Michea ire debemus in Ramoth Galaad ad they shall be delivered into your hand. bellandum an quiescere cui ille respondit ascendite cuncta

Biblia Sacra Vulgata

enim prospera euenient et tradentur hostes in manus uestras

18:15 And the king said to him, How many times shall I 18:15 dixitque rex iterum atque iterum te adiuro ut non mihi

adjure thee that thou say nothing but the truth to me in the name of the LORD?

loquaris nisi quod uerum est in nomine Domini 18:16 at ille

18:16 Then he said, I did see all Israel scattered upon the mountains, as sheep that have no shepherd: and the LORD said, These have no master; let them return therefore every man to his house in peace.

ait uidi uniuersum Israhel dispersum in montibus sicut oues absque pastore et dixit Dominus non habent isti dominos re-

18:17 And the king of Israel said to Jehoshaphat, Did I uertatur unusquisque ad domum suam in pace 18:17 et ait

not tell thee that he would not prophesy good unto me, but evil?

rex Israhel ad Iosaphat nonne dixi tibi quod non prophetaret

18:18 Again he said, Therefore hear the word of the iste mihi quicquam boni sed ea quae mala sunt 18:18 at ille

LORD; I saw the LORD sitting upon his throne, and all the host of heaven standing on his right hand and on his

left. idcirco ait audite uerbum Domini uidi Dominum sedentem

in solio suo et omnem exercitum caeli adsistentem ei a dext-

18:19 And the LORD said, Who shall entice Ahab king of ris et sinistris 18:19 et dixit Dominus quis decipiet Ahab re-

Israel, that he may go up and fall at Ramothgilead? And one spake saying after this manner, and another saying af-

ter that manner. gem Israhel ut ascendat et corruat in Ramoth Galaad cumque 18:20 Then there came out a spirit, and stood before the diceret unus hoc modo et alter alio 18:20 processit spiritus

LORD, and said, I will entice him. And the LORD said unto him, Wherewith?

et stetit coram Domino et ait ego decipiam eum cui Dominus

18:21 And he said, I will go out, and be a lying spirit in in quo inquit decipies 18:21 at ille respondit egrediar et ero

the mouth of all his prophets. And the Lord said, Thou shalt entice him, and thou shalt also prevail: go out, and

do even so. spiritus mendax in ore omnium prophetarum eius dixitque 18:22 Now therefore, behold, the LORD hath put a lying Dominus decipies et praeualebis egredere et fac ita 18:22

spirit in the mouth of these thy prophets, and the LORD hath spoken evil against thee.

nunc igitur ecce dedit Dominus spiritum mendacii in ore omnium prophetarum tuorum et Dominus locutus est de te mala

18:23 Then Zedekiah the son of Chenaanah came near, 18:23 accessit autem Sedecias filius Chanana et percussit Mi-

and smote Micaiah upon the cheek, and said, Which way went the Spirit of the LORD from me to speak unto thee?

cheae maxillam et ait per quam uiam transiuit spiritus Do-

18:24 And Micaiah said, Behold, thou shalt see on that day mini a me ut loqueretur tibi 18:24 dixitque Micheas tu ipse

when thou shalt go into an inner chamber to hide thyself.

uidebis in die illo quando ingressus fueris cubiculum de cu-

18:25 Then the king of Israel said, Take ye Micaiah, and biculo ut abscondaris 18:25 praecepit autem rex Israhel di-

carry him back to Amon the governor of the city, and to Joash the king’s son;

cens tollite Micheam et ducite eum ad Amon principem ci-

18:26 And say, Thus saith the king, Put this fellow in the uitatis et ad Ioas filium Ammelech 18:26 et dicetis haec di-

prison, and feed him with bread of affliction and with water of affliction, until I return in peace.

cit rex mittite hunc in carcerem et date ei panis modicum et

18:27 And Micaiah said, If thou certainly return in peace, aquae pauxillum donec reuertar in pace 18:27 dixitque Mithen hath not the LORD spoken by me. And he said, Hearken, all ye people.

cheas si reuersus fueris in pace non est locutus Dominus in

18:28 So the king of Israel and Jehoshaphat the king of me et ait audite populi omnes 18:28 igitur ascenderunt rex

Judah went up to Ramothgilead.

18:29 And the king of Israel said unto Jehoshaphat, I will Israhel et Iosaphat rex Iuda in Ramoth Galaad 18:29 dixit-

disguise myself, and I will go to the battle; but put thou on thy robes. So the king of Israel disguised himself; and

they went to the battle. que rex Israhel ad Iosaphat mutabo habitum et sic ad pugnandum uadam tu autem induere uestibus tuis mutatoque

Liber II Paralipomenon

rex Israhel habitu uenit ad bellum 18:30 rex autem Syriae 18:30 Now the king of Syria had commanded the captains

of the chariots that were with him, saying, Fight ye not with small or great, save only with the king of Israel.

praeceperat ducibus equitatus sui dicens ne pugnetis contra minimum aut contra maximum nisi contra solum regem Israhel 18:31 itaque cum uidissent principes equitatus Iosaphat 18:31 And it came to pass, when the captains of the char-

iots saw Jehoshaphat, that they said, It is the king of Israel. Therefore they compassed about him to fight: but Jehoshaphat cried out, and the LORD helped him; and God moved them to depart from him.

dixerunt rex Israhel iste est et circumdederunt eum dimicantes at ille clamauit ad Dominum et auxiliatus est ei atque auertit eos ab illo 18:32 cum enim uidissent duces equita- 18:32 For it came to pass, that, when the captains of the

chariots perceived that it was not the king of Israel, they turned back again from pursuing him.

tus quod non esset rex Israhel reliquerunt eum 18:33 accidit

18:33 And a certain man drew a bow at a venture, and smote the king of Israel between the joints of the harness: therefore he said to his chariot man, Turn thine hand, that thou mayest carry me out of the host; for I am wounded.

autem ut unus e populo sagittam in incertum iaceret et percuteret regem Israhel inter ceruicem et scapulas at ille aurigae suo ait conuerte manum tuam et educ me de acie quia uul-

neratus sum 18:34 et finita est pugna in die illo porro rex 18:34 And the battle increased that day: howbeit the king of Israel stayed himself up in his chariot against the Syrians until the even: and about the time of the sun going

Israhel stabat in curru suo contra Syros usque ad uesperam down he died. et mortuus est occidente sole

19:1 reuersus est autem Iosaphat rex Iuda domum suam 19:1 And Jehoshaphat the king of Judah returned to his

house in peace to Jerusalem.

pacifice in Hierusalem 19:2 cui occurrit Hieu filius Anani 19:2 And Jehu the son of Hanani the seer went out to meet

him, and said to king Jehoshaphat, Shouldest thou help the ungodly, and love them that hate the LORD? therefore is

uidens et ait ad eum impio praebes auxilium et his qui ode- wrath upon thee from before the LORD. runt Dominum amicitia iungeris et idcirco iram quidem Do-

mini merebaris 19:3 sed bona opera inuenta sunt in te eo 19:3 Nevertheless there are good things found in thee, in that thou hast taken away the groves out of the land, and hast prepared thine heart to seek God.

quod abstuleris lucos de terra Iuda et praeparaueris cor tuum ut requireres Dominum 19:4 habitauit ergo Iosaphat in Hie- 19:4 And Jehoshaphat dwelt at Jerusalem: and he went

out again through the people from Beersheba to mount Ephraim, and brought them back unto the LORD God of

rusalem rursumque egressus est ad populum de Bersabee their fathers. usque ad montem Ephraim et reuocauit eos ad Dominum Deum patrum suorum 19:5 constituitque iudices terrae in 19:5 And he set judges in the land throughout all the

fenced cities of Judah, city by city,

cunctis ciuitatibus Iuda munitis per singula loca 19:6 et 19:6 And said to the judges, Take heed what ye do: for ye

judge not for man, but for the LORD, who is with you in the judgment.

praecipiens iudicibus uidete ait quid faciatis non enim hominis exercetis iudicium sed Domini et quodcumque iudicaueritis in uos redundabit 19:7 sit timor Domini uobiscum 19:7 Wherefore now let the fear of the LORD be upon

you; take heed and do it: for there is no iniquity with the LORD our God, nor respect of persons, nor taking of gifts.

et cum diligentia cuncta facite non est enim apud Dominum Deum nostrum iniquitas nec personarum acceptio nec cupido munerum 19:8 in Hierusalem quoque constituit Iosa- 19:8 Moreover in Jerusalem did Jehoshaphat set of the

Levites, and of the priests, and of the chief of the fathers of Israel, for the judgment of the LORD, and for contro-

phat Leuitas et sacerdotes et principes familiarum ex Israhel versies, when they returned to Jerusalem. ut iudicium et causam Domini iudicarent habitatoribus eius

Biblia Sacra Vulgata

19:9 And he charged them, saying, Thus shall ye do in the 19:9 praecepitque eis dicens sic agetis in timore Dei fideliter

fear of the LORD, faithfully, and with a perfect heart.

19:10 And what cause soever shall come to you of your et corde perfecto 19:10 omnem causam quae uenerit ad uos

brethren that dwell in your cities, between blood and blood, between law and commandment, statutes and judgments, ye shall even warn them that they trespass not against the LORD, and so wrath come upon you, and upon your brethren: this do, and ye shall not trespass.

fratrum uestrorum qui habitant in urbibus suis inter cognationem et cognationem ubicumque quaestio est de lege de mandato de caerimoniis de iustificationibus ostendite eis ut non peccent in Dominum et ne ueniat ira super uos et super

19:11 And, behold, Amariah the chief priest is over you fratres uestros sic ergo agetis et non peccabitis 19:11 Ama-

in all matters of the LORD; and Zebadiah the son of Ishmael, the ruler of the house of Judah, for all the king’s matters: also the Levites shall be officers before you. Deal courageously, and the LORD shall be with the good.

rias autem sacerdos et pontifex uester in his quae ad Dominum pertinent praesidebit porro Zabadias filius Ismahel qui est dux in domo Iuda super ea opera erit quae ad regis officium pertinent habetisque magistros Leuitas coram uobis confortamini et agite diligenter et erit Dominus cum bonis

20:1 It came to pass after this also, that the children of 20:1 post haec congregati sunt filii Moab et filii Ammon Moab, and the children of Ammon, and with them other beside the Ammonites, came against Jehoshaphat to battle.

et cum eis de Ammanitis ad Iosaphat ut pugnarent contra

20:2 Then there came some that told Jehoshaphat, saying, eum 20:2 ueneruntque nuntii et indicauerunt Iosaphat di-

There cometh a great multitude against thee from beyond the sea on this side Syria; and, behold, they be in Haza-

zontamar, which is Engedi. centes uenit contra te multitudo magna de his locis quae

trans mare sunt et de Syria et ecce consistunt in Asasont-

20:3 And Jehoshaphat feared, and set himself to seek the hamar quae est Engaddi 20:3 Iosaphat autem timore perter-

LORD, and proclaimed a fast throughout all Judah.

ritus totum se contulit ad rogandum Dominum et praedicauit

20:4 And Judah gathered themselves together, to ask help ieiunium uniuerso Iuda 20:4 congregatusque Iudas ad pre-

of the LORD: even out of all the cities of Judah they came to seek the LORD.

candum Dominum sed et omnes de urbibus suis uenerunt ad

20:5 And Jehoshaphat stood in the congregation of Judah obsecrandum eum 20:5 cumque stetisset Iosaphat in medio

and Jerusalem, in the house of the LORD, before the new court,

coetu Iudae et Hierusalem in domo Domini ante atrium no-

20:6 And said, O LORD God of our fathers, art not thou uum 20:6 ait Domine Deus patrum nostrorum tu es Deus God in heaven? and rulest not thou over all the kingdoms of the heathen? and in thine hand is there not power and

might, so that none is able to withstand thee? in caelo et dominaris cunctis regnis gentium in manu tua est 20:7 Art not thou our God, who didst drive out the inhabi- fortitudo et potentia nec quisquam tibi potest resistere 20:7

tants of this land before thy people Israel, and gavest it to the seed of Abraham thy friend for ever?

nonne tu Deus noster interfecisti omnes habitatores terrae huius coram populo tuo Israhel et dedisti eam semini Abra-

20:8 And they dwelt therein, and have built thee a sanctu- ham amici tui in sempiternum 20:8 habitaueruntque in ea et

ary therein for thy name, saying,

20:9 If, when evil cometh upon us, as the sword, judgment, extruxerunt in illa sanctuarium nomini tuo dicentes 20:9 si

or pestilence, or famine, we stand before this house, and in thy presence, (for thy name is in this house,) and cry

unto thee in our affliction, then thou wilt hear and help. inruerint super nos mala gladius iudicii pestilentia et fames stabimus coram domo hac in conspectu tuo in qua inuocatum est nomen tuum et clamabimus ad te in tribulationibus

Liber II Paralipomenon

nostris et exaudies saluosque facies 20:10 nunc igitur ecce 20:10 And now, behold, the children of Ammon and Moab

and mount Seir, whom thou wouldest not let Israel invade, when they came out of the land of Egypt, but they turned

filii Ammon et Moab et mons Seir per quos non concessisti from them, and destroyed them not; Israheli ut transirent quando egrediebantur de Aegypto sed declinauerunt ab eis et non interfecerunt illos 20:11 e con- 20:11 Behold, I say, how they reward us, to come to cast us

out of thy possession, which thou hast given us to inherit.

trario agunt et nituntur eicere nos de possessione quam tradidisti nobis 20:12 Deus noster ergo non iudicabis eos in 20:12 O our God, wilt thou not judge them? for we have

no might against this great company that cometh against us; neither know we what to do: but our eyes are upon

nobis quidem non tanta est fortitudo ut possimus huic mul- thee. titudini resistere quae inruit super nos sed cum ignoremus quid agere debeamus hoc solum habemus residui ut oculos nostros dirigamus ad te 20:13 omnis uero Iuda stabat coram 20:13 And all Judah stood before the LORD, with their

little ones, their wives, and their children.

Domino cum paruulis et uxoribus et liberis suis 20:14 erat 20:14 Then upon Jahaziel the son of Zechariah, the son of

Benaiah, the son of Jeiel, the son of Mattaniah, a Levite of the sons of Asaph, came the Spirit of the LORD in the

autem Hiazihel filius Zacchariae filii Banaiae filii Hiehihel midst of the congregation; filii Mathaniae Leuites de filiis Asaph super quem factus est spiritus Domini in medio turbae 20:15 et ait adtendite om- 20:15 And he said, Hearken ye, all Judah, and ye inhabi-

tants of Jerusalem, and thou king Jehoshaphat, Thus saith the LORD unto you, Be not afraid nor dismayed by reason of this great multitude; for the battle is not yours, but God’s.

nis Iuda et qui habitatis Hierusalem et tu rex Iosaphat haec dicit Dominus uobis nolite timere nec paueatis hanc multitudinem non est enim uestra pugna sed Dei 20:16 cras de- 20:16 To morrow go ye down against them: behold, they

come up by the cliff of Ziz; and ye shall find them at the end of the brook, before the wilderness of Jeruel.

scendetis contra eos ascensuri enim sunt per cliuum nomine Sis et inuenietis illos in summitate torrentis qui est contra solitudinem Hieruhel 20:17 non eritis uos qui dimicabitis sed 20:17 Ye shall not need to fight in this battle: set your-

selves, stand ye still, and see the salvation of the LORD with you, O Judah and Jerusalem: fear not, nor be dismayed; to morrow go out against them: for the LORD will be with you.

tantummodo confidenter state et uidebitis auxilium Domini super uos o Iuda et Hierusalem nolite timere nec paueatis cras egredimini contra eos et Dominus erit uobiscum 20:18 20:18 And Jehoshaphat bowed his head with his face to

the ground: and all Judah and the inhabitants of Jerusalem fell before the LORD, worshipping the LORD.

Iosaphat ergo et Iuda et omnes habitatores Hierusalem ceciderunt proni in terram coram Domino et adorauerunt eum 20:19 porro Leuitae de filiis Caath et de filiis Core laudabant 20:19 And the Levites, of the children of the Kohathites,

and of the children of the Korhites, stood up to praise the LORD God of Israel with a loud voice on high.

Dominum Deum Israhel uoce magna in excelsum 20:20

20:20 And they rose early in the morning, and went forth into the wilderness of Tekoa: and as they went forth, Jehoshaphat stood and said, Hear me, O Judah, and ye inhabitants of Jerusalem; Believe in the LORD your God, so shall ye be established; believe his prophets, so shall ye prosper.

cumque mane surrexissent egressi sunt per desertum Thecuae profectisque eis stans Iosaphat in medio eorum dixit audite me Iuda et omnes habitatores Hierusalem credite in Domino Deo uestro et securi eritis credite prophetis eius et cuncta euenient prospera 20:21 deditque consilium populo 20:21 And when he had consulted with the people, he ap-

pointed singers unto the LORD, and that should praise the beauty of holiness, as they went out before the army, and

et statuit cantores Domini ut laudarent eum in turmis suis to say, Praise the LORD; for his mercy endureth for ever.

Biblia Sacra Vulgata

et antecederent exercitum ac uoce consona dicerent confite-

20:22 And when they began to sing and to praise, the mini Domino quoniam in aeternum misericordia eius 20:22

LORD set ambushments against the children of Ammon, Moab, and mount Seir, which were come against Judah;

and they were smitten. cumque coepissent laudes canere uertit Dominus insidias eorum in semet ipsos filiorum scilicet Ammon et Moab et montis Seir qui egressi fuerant ut pugnarent contra Iudam et

20:23 For the children of Ammon and Moab stood up percussi sunt 20:23 namque filii Ammon et Moab consur-

against the inhabitants of mount Seir, utterly to slay and destroy them: and when they had made an end of the in-

habitants of Seir, every one helped to destroy another. rexerunt aduersum habitatores montis Seir ut interficerent et

delerent eos cumque hoc opere perpetrassent etiam in semet

20:24 And when Judah came toward the watch tower in ipsos uersi mutuis concidere uulneribus 20:24 porro Iudas

the wilderness, they looked unto the multitude, and, behold, they were dead bodies fallen to the earth, and none

escaped. cum uenisset ad speculam quae respicit solitudinem uidit

procul omnem late regionem plenam cadaueribus nec supe-

20:25 And when Jehoshaphat and his people came to take resse quemquam qui necem potuisset euadere 20:25 uenit

away the spoil of them, they found among them in abundance both riches with the dead bodies, and precious jewels, which they stripped off for themselves, more than they could carry away: and they were three days in gathering of the spoil, it was so much.

ergo Iosaphat et omnis populus cum eo ad detrahenda spolia mortuorum inueneruntque inter cadauera uariam supellectilem uestes quoque et uasa pretiosissima et diripuerunt ita ut omnia portare non possent nec per tres dies spolia auferre

20:26 And on the fourth day they assembled themselves in pro praedae magnitudine 20:26 die autem quarto congrethe valley of Berachah; for there they blessed the LORD: therefore the name of the same place was called, The val-

ley of Berachah, unto this day. gati sunt in ualle Benedictionis etenim quoniam ibi bene-

dixerant Domino uocauerunt locum illum uallis Benedictio-

20:27 Then they returned, every man of Judah and nis usque in praesentem diem 20:27 reuersusque est omnis

Jerusalem, and Jehoshaphat in the forefront of them, to go again to Jerusalem with joy; for the LORD had made

them to rejoice over their enemies. uir Iuda et habitatores Hierusalem et Iosaphat ante eos in

Hierusalem cum laetitia magna eo quod dedisset eis Domi-

20:28 And they came to Jerusalem with psalteries and nus gaudium de inimicis suis 20:28 ingressique sunt Hieru-

harps and trumpets unto the house of the LORD.

salem cum psalteriis et citharis et tubis in domum Domini

20:29 And the fear of God was on all the kingdoms of 20:29 inruit autem pauor Domini super uniuersa regna terra-

those countries, when they had heard that the LORD fought against the enemies of Israel.

rum cum audissent quod pugnasset Dominus contra inimicos

20:30 So the realm of Jehoshaphat was quiet: for his God Israhel 20:30 quieuitque regnum Iosaphat et praebuit ei De-

gave him rest round about.

20:31 And Jehoshaphat reigned over Judah: he was thirty us pacem per circuitum 20:31 regnauit igitur Iosaphat super

and five years old when he began to reign, and he reigned twenty and five years in Jerusalem. And his mother’s

name was Azubah the daughter of Shilhi. Iudam et erat triginta quinque annorum cum regnare coepis-

set uiginti autem et quinque annis regnauit in Hierusalem

20:32 And he walked in the way of Asa his father, and nomen matris eius Azuba filia Selachi 20:32 et ambulauit

departed not from it, doing that which was right in the sight of the LORD.

in uia patris sui Asa nec declinauit ab ea faciens quae placita

20:33 Howbeit the high places were not taken away: for as erant coram Domino 20:33 uerumtamen excelsa non absyet the people had not prepared their hearts unto the God of their fathers.

Liber II Paralipomenon

tulit et adhuc populus non direxerat cor suum ad Dominum Deum patrum suorum 20:34 reliqua autem gestorum Iosa- 20:34 Now the rest of the acts of Jehoshaphat, first and

last, behold, they are written in the book of Jehu the son of Hanani, who is mentioned in the book of the kings of

phat priorum et nouissimorum scripta sunt in uerbis Hieu Israel. filii Anani quae digessit in libro regum Israhel 20:35 post 20:35 And after this did Jehoshaphat king of Judah

join himself with Ahaziah king of Israel, who did very wickedly:

haec iniit amicitias Iosaphat rex Iuda cum Ochozia rege Israhel cuius opera fuerunt impiissima 20:36 et particeps fuit 20:36 And he joined himself with him to make ships to go

to Tarshish: and they made the ships in Eziongaber.

ut facerent naues quae irent in Tharsis feceruntque classem in Asiongaber 20:37 prophetauit autem Eliezer filius Do- 20:37 Then Eliezer the son of Dodavah of Mareshah

prophesied against Jehoshaphat, saying, Because thou hast joined thyself with Ahaziah, the LORD hath broken thy works. And the ships were broken, that they were not able to go to Tarshish.

doau de Maresa ad Iosaphat dicens quia habuisti foedus cum Ochozia percussit Dominus opera tua contritaeque sunt naues nec potuerunt ire in Tharsis

21:1 dormiuit autem Iosaphat cum patribus suis et sepul- 21:1 Now Jehoshaphat slept with his fathers, and was

buried with his fathers in the city of David. And Jehoram his son reigned in his stead.

tus est cum eis in ciuitate Dauid regnauitque Ioram filius eius pro eo 21:2 qui habuit fratres filios Iosaphat Azariam 21:2 And he had brethren the sons of Jehoshaphat,

Azariah, and Jehiel, and Zechariah, and Azariah, and Michael, and Shephatiah: all these were the sons of Je-

et Hiahihel et Zacchariam et Azariam et Michahel et Sapha- hoshaphat king of Israel. tiam omnes hii filii Iosaphat regis Israhel 21:3 deditque eis 21:3 And their father gave them great gifts of silver, and of

gold, and of precious things, with fenced cities in Judah: but the kingdom gave he to Jehoram; because he was the

pater suus multa munera argenti et auri et pensitationes cum firstborn. ciuitatibus munitissimis in Iuda regnum autem tradidit Ioram eo quod esset primogenitus 21:4 surrexit ergo Ioram 21:4 Now when Jehoram was risen up to the kingdom

of his father, he strengthened himself, and slew all his brethren with the sword, and divers also of the princes of

super regnum patris sui cumque se confirmasset occidit om- Israel. nes fratres suos gladio et quosdam de principibus Israhel 21:5 triginta duo annorum erat Ioram cum regnare coepis- 21:5 Jehoram was thirty and two years old when he began

to reign, and he reigned eight years in Jerusalem.

set et octo annis regnauit in Hierusalem 21:6 ambulauitque 21:6 And he walked in the way of the kings of Israel, like

as did the house of Ahab: for he had the daughter of Ahab to wife: and he wrought that which was evil in the eyes of

in uiis regum Israhel sicut egerat domus Ahab filia quippe the LORD. Ahab erat uxor eius et fecit malum in conspectu Domini 21:7 noluit autem Dominus disperdere domum Dauid pro- 21:7 Howbeit the LORD would not destroy the house of

David, because of the covenant that he had made with David, and as he promised to give a light to him and to

pter pactum quod inierat cum eo et quia promiserat ut daret his sons for ever. illi lucernam et filiis eius omni tempore 21:8 in diebus il- 21:8 In his days the Edomites revolted from under the do-

minion of Judah, and made themselves a king.

lis rebellauit Edom ne esset subditus Iudae et constituit sibi regem 21:9 cumque transisset Ioram cum principibus suis 21:9 Then Jehoram went forth with his princes, and all his

chariots with him: and he rose up by night, and smote the Edomites which compassed him in, and the captains of the

et cuncto equitatu qui erat secum surrexit nocte et percussit chariots. Edom qui se circumdederat et omnes duces equitatus eius 21:10 attamen rebellauit Edom ne esset sub dicione Iuda us- 21:10 So the Edomites revolted from under the hand of

Judah unto this day. The same time also did Libnah revolt from under his hand; because he had forsaken the LORD God of his fathers.

Biblia Sacra Vulgata

que ad hanc diem eo tempore et Lobna recessit ne esset sub manu illius dereliquerat enim Dominum Deum patrum suo-

21:11 Moreover he made high places in the mountains of rum 21:11 insuper et excelsa fabricatus est in urbibus Iuda

Judah and caused the inhabitants of Jerusalem to commit fornication, and compelled Judah thereto.

et fornicari fecit habitatores Hierusalem et praeuaricari Iu-

21:12 And there came a writing to him from Elijah the dam 21:12 adlatae sunt autem ei litterae ab Helia propheta

prophet, saying, Thus saith the LORD God of David thy father, Because thou hast not walked in the ways of Jehoshaphat thy father, nor in the ways of Asa king of Judah, in quibus scriptum erat haec dicit Dominus Deus Dauid pa-

tris tui quoniam non ambulasti in uiis Iosaphat patris tui et

21:13 But hast walked in the way of the kings of Israel, and in uiis Asa regis Iuda 21:13 sed incessisti per iter regum hast made Judah and the inhabitants of Jerusalem to go a whoring, like to the whoredoms of the house of Ahab, and also hast slain thy brethren of thy father’s house, which were better than thyself:

Israhel et fornicari fecisti Iudam et habitatores Hierusalem imitatus fornicationem domus Ahab insuper et fratres tuos

21:14 Behold, with a great plague will the LORD smite domum patris tui meliores te occidisti 21:14 ecce Dominus

thy people, and thy children, and thy wives, and all thy goods:

percutiet te plaga magna cum populo tuo et filiis et uxoribus

21:15 And thou shalt have great sickness by disease of thy tuis uniuersaque substantia tua 21:15 tu autem aegrotabis

bowels, until thy bowels fall out by reason of the sickness day by day.

pessimo languore uteri donec egrediantur uitalia tua paula-

21:16 Moreover the LORD stirred up against Jehoram the tim per dies singulos 21:16 suscitauit ergo Dominus contra

spirit of the Philistines, and of the Arabians, that were near the Ethiopians:

Ioram spiritum Philisthinorum et Arabum qui confines sunt

21:17 And they came up into Judah, and brake into it, and Aethiopibus 21:17 et ascenderunt in terram Iuda et uastaue-

carried away all the substance that was found in the king’s house, and his sons also, and his wives; so that there was never a son left him, save Jehoahaz, the youngest of his sons.

runt eam diripueruntque cunctam substantiam quae inuenta est in domo regis insuper et filios eius et uxores nec remansit

21:18 And after all this the LORD smote him in his bowels ei filius nisi Ioachaz qui minimus natu erat 21:18 et super

with an incurable disease.

haec omnia percussit eum Dominus alui languore insanabili

21:19 And it came to pass, that in process of time, after 21:19 cumque diei succederet dies et temporum spatia uol-

the end of two years, his bowels fell out by reason of his sickness: so he died of sore diseases. And his people made

no burning for him, like the burning of his fathers. uerentur duorum annorum expletus est circulus et sic longa consumptus tabe ita ut egereret etiam uiscera sua languore pariter et uita caruit mortuusque est in infirmitate pessima et non fecit ei populus secundum morem conbustionis exequias

21:20 Thirty and two years old was he when he began to sicut fecerat maioribus eius 21:20 triginta duum annorum

reign, and he reigned in Jerusalem eight years, and departed without being desired. Howbeit they buried him in

the city of David, but not in the sepulchres of the kings. fuit cum regnare coepisset et octo annis regnauit in Hierusalem ambulauitque non recte et sepelierunt eum in ciuitate Dauid uerumtamen non in sepulchro regum

22:1 And the inhabitants of Jerusalem made Ahaziah his 22:1 constituerunt autem habitatores Hierusalem Ocho-

youngest son king in his stead: for the band of men that came with the Arabians to the camp had slain all the eldest.

So Ahaziah the son of Jehoram king of Judah reigned. ziam filium eius minimum regem pro eo omnes enim maiores natu qui ante eum fuerant interfecerant latrones Arabum

Liber II Paralipomenon

qui inruerant in castra regnauitque Ochozias filius Ioram regis Iuda 22:2 filius quadraginta duo annorum erat Ochozias 22:2 Forty and two years old was Ahaziah when he be-

gan to reign, and he reigned one year in Jerusalem. His mother’s name also was Athaliah the daughter of Omri.

cum regnare coepisset et uno anno regnauit in Hierusalem nomen matris eius Otholia filia Amri 22:3 sed et ipse in- 22:3 He also walked in the ways of the house of Ahab: for

his mother was his counsellor to do wickedly.

gressus est per uias domus Ahab mater enim eius inpulit eum ut impie ageret 22:4 fecit igitur malum in conspectu 22:4 Wherefore he did evil in the sight of the LORD like

the house of Ahab: for they were his counsellors after the death of his father to his destruction.

Domini sicut domus Ahab ipsi enim fuerunt ei consiliarii post mortem patris sui in interitum eius 22:5 ambulauitque 22:5 He walked also after their counsel, and went with Je-

horam the son of Ahab king of Israel to war against Hazael king of Syria at Ramothgilead: and the Syrians smote Jo-

in consiliis eorum et perrexit cum Ioram filio Ahab rege Isra- ram. hel in bellum contra Azahel regem Syriae in Ramoth Galaad uulneraueruntque Syri Ioram 22:6 qui reuersus est ut cu- 22:6 And he returned to be healed in Jezreel because of the

wounds which were given him at Ramah, when he fought with Hazael king of Syria. And Azariah the son of Jehoram king of Judah went down to see Jehoram the son of Ahab at Jezreel, because he was sick.

raretur in Hiezrahel multas enim plagas acceperat in supradicto certamine igitur Azarias filius Ioram rex Iuda descendit ut inuiseret Ioram filium Ahab in Hiezrahel aegrotantem 22:7 uoluntatis quippe fuit Dei aduersum Ochoziam ut ueni- 22:7 And the destruction of Ahaziah was of God by com-

ing to Joram: for when he was come, he went out with Jehoram against Jehu the son of Nimshi, whom the LORD

ret ad Ioram et cum uenisset egrederetur cum eo aduersum had anointed to cut off the house of Ahab. Hieu filium Namsi quem unxit Dominus ut deleret domum Ahab 22:8 cum ergo subuerteret Hieu domum Ahab inuenit 22:8 And it came to pass, that, when Jehu was executing

judgment upon the house of Ahab, and found the princes of Judah, and the sons of the brethren of Ahaziah, that

principes Iuda et filios fratrum Ochoziae qui ministrabant ei ministered to Ahaziah, he slew them. et interfecit illos 22:9 ipsumque perquirens Ochoziam con- 22:9 And he sought Ahaziah: and they caught him, (for he

was hid in Samaria,) and brought him to Jehu: and when they had slain him, they buried him: Because, said they, he is the son of Jehoshaphat, who sought the LORD with all his heart. So the house of Ahaziah had no power to keep still the kingdom.

prehendit latentem in Samaria adductumque ad se occidit et sepelierunt eum eo quod esset filius Iosaphat qui quaesierat Dominum in toto corde suo nec erat ultra spes aliqua ut de stirpe regnaret Ochoziae 22:10 siquidem Otholia mater 22:10 But when Athaliah the mother of Ahaziah saw that

her son was dead, she arose and destroyed all the seed royal of the house of Judah.

eius uidens quod mortuus esset filius suus surrexit et interfecit omnem stirpem regiam domus Ioram 22:11 porro Iosa- 22:11 But Jehoshabeath, the daughter of the king, took

Joash the son of Ahaziah, and stole him from among the king’s sons that were slain, and put him and his nurse in a bedchamber. So Jehoshabeath, the daughter of king Jehoram, the wife of Jehoiada the priest, (for she was the sister of Ahaziah,) hid him from Athaliah, so that she slew him not.

beth filia regis tulit Ioas filium Ochoziae et furata est eum de medio filiorum regis cum interficerentur absconditque cum nutrice sua in cubiculo lectulorum Iosabeth autem quae absconderat eum erat filia regis Ioram uxor Ioiadae pontificis soror Ochoziae et idcirco Otholia non interfecit eum 22:12 22:12 And he was with them hid in the house of God six

years: and Athaliah reigned over the land.

fuit ergo cum eis in domo Dei absconditus sex annis quibus regnauit Otholia super terram

Biblia Sacra Vulgata

23:1 And in the seventh year Jehoiada strengthened him- 23:1 anno autem septimo confortatus Ioiadae adsumpsit self, and took the captains of hundreds, Azariah the son of Jeroham, and Ishmael the son of Jehohanan, and Azariah the son of Obed, and Maaseiah the son of Adaiah, and Elishaphat the son of Zichri, into covenant with him.

centuriones Azariam uidelicet filium Hieroam et Ismahel filium Iohanan Azariam quoque filium Oded et Maasiam filium Adaiae et Elisaphat filium Zechri et iniit cum eis foe-

23:2 And they went about in Judah, and gathered the dus 23:2 qui circumeuntes Iudam congregauerunt Leuitas

Levites out of all the cities of Judah, and the chief of the fathers of Israel, and they came to Jerusalem.

de cunctis urbibus Iuda et principes familiarum Israhel ue-

23:3 And all the congregation made a covenant with the neruntque in Hierusalem 23:3 iniit igitur omnis multitudo

king in the house of God. And he said unto them, Behold, the king’s son shall reign, as the LORD hath said of the

sons of David. pactum in domo Domini cum rege dixitque ad eos Ioiadae

ecce filius regis regnabit sicut locutus est Dominus super fi-

23:4 This is the thing that ye shall do; A third part of you lios Dauid 23:4 iste est ergo sermo quem facietis 23:5 ter-

entering on the sabbath, of the priests and of the Levites, shall be porters of the doors; 23:5 And a third part shall be at the king’s house; and a third part at the gate of the foundation: and all the people shall be in the courts of the house of the LORD.

tia pars uestrum qui ueniunt ad sabbatum sacerdotum et Leuitarum et ianitorum erit in portis tertia uero pars ad domum regis et tertia in porta quae appellatur Fundamenti omne

23:6 But let none come into the house of the LORD, save uero reliquum uulgus sit in atriis domus Domini 23:6 nec

the priests, and they that minister of the Levites; they shall go in, for they are holy: but all the people shall keep the

watch of the LORD. quisquam alius ingrediatur domum Domini nisi sacerdotes et qui ministrant de Leuitis ipsi tantummodo ingrediantur quia sanctificati sunt et omne reliquum uulgus obseruet cu-

23:7 And the Levites shall compass the king round about, stodias Domini 23:7 Leuitae autem circumdent regem haevery man with his weapons in his hand; and whosoever else cometh into the house, he shall be put to death: but be ye with the king when he cometh in, and when he goeth out.

bentes singuli arma sua et si quis alius ingressus fuerit templum interficiatur sintque cum rege et intrante et egrediente

23:8 So the Levites and all Judah did according to all 23:8 fecerunt igitur Leuitae et uniuersus Iuda iuxta omnia

things that Jehoiada the priest had commanded, and took every man his men that were to come in on the sabbath, with them that were to go out on the sabbath: for Jehoiada the priest dismissed not the courses.

quae praeceperat Ioiadae pontifex et adsumpserunt singuli uiros qui sub se erant et ueniebant per ordinem sabbati cum his qui iam impleuerant sabbatum et egressuri erant siquidem Ioiadae pontifex non dimiserat abire turmas quae sibi

23:9 Moreover Jehoiada the priest delivered to the cap- per singulas ebdomadas succedere consueuerant 23:9 de-

tains of hundreds spears, and bucklers, and shields, that had been king David’s, which were in the house of God.

ditque Ioiadae sacerdos centurionibus lanceas clypeosque et peltas regis Dauid quas consecrauerat in domo Domini

23:10 And he set all the people, every man having his 23:10 constituitque omnem populum tenentium pugiones a

weapon in his hand, from the right side of the temple to the left side of the temple, along by the altar and the tem-

ple, by the king round about. parte templi dextra usque ad partem templi sinistram coram 23:11 Then they brought out the king’s son, and put upon altari et templo per circuitum regis 23:11 et eduxerunt fi-

him the crown, and gave him the testimony, and made him king. And Jehoiada and his sons anointed him, and said,

God save the king. lium regis et inposuerunt ei diadema dederuntque in manu eius tenendam legem et constituerunt eum regem unxit quo-

Liber II Paralipomenon

que illum Ioiadae pontifex et filii eius inprecatique sunt atque dixerunt uiuat rex 23:12 quod cum audisset Otholia uo- 23:12 Now when Athaliah heard the noise of the people

running and praising the king, she came to the people into the house of the LORD:

cem scilicet currentium atque laudantium regem ingressa est ad populum in templum Domini 23:13 cumque uidisset re- 23:13 And she looked, and, behold, the king stood at his

pillar at the entering in, and the princes and the trumpets by the king: and all the people of the land rejoiced, and sounded with trumpets, also the singers with instruments of musick, and such as taught to sing praise. Then Athaliah rent her clothes, and said, Treason, Treason.

gem stantem super gradum in introitu et principes turmasque circa eum omnem quoque populum terrae gaudentem atque clangentem tubis et diuersi generis organis concinentem uocemque laudantium scidit uestimenta sua et ait insidiae insidiae 23:14 egressus autem Ioiadae pontifex ad cen- 23:14 Then Jehoiada the priest brought out the captains of

hundreds that were set over the host, and said unto them, Have her forth of the ranges: and whoso followeth her, let him be slain with the sword. For the priest said, Slay her not in the house of the LORD.

turiones et principes exercitus dixit eis educite illam extra septa templi et interficiatur foris gladio praecepitque sacerdos ne occideretur in domo Domini 23:15 et inposuerunt 23:15 So they laid hands on her; and when she was come

to the entering of the horse gate by the king’s house, they slew her there.

ceruicibus eius manus cumque intrasset portam Equorum domus regis interfecerunt eam ibi 23:16 pepigit autem Io- 23:16 And Jehoiada made a covenant between him, and

between all the people, and between the king, that they should be the LORD’s people.

iadae foedus inter se uniuersumque populum et regem ut esset populus Domini 23:17 itaque ingressus est omnis po- 23:17 Then all the people went to the house of Baal, and

brake it down, and brake his altars and his images in pieces, and slew Mattan the priest of Baal before the al-

pulus domum Baal et destruxerunt eam et altaria ac simu- tars. lacra illius confregerunt Matthan quoque sacerdotem Baal interfecerunt ante aras 23:18 constituit autem Ioiadae prae- 23:18 Also Jehoiada appointed the offices of the house of

the LORD by the hand of the priests the Levites, whom David had distributed in the house of the LORD, to offer the burnt offerings of the LORD, as it is written in the law of Moses, with rejoicing and with singing, as it was ordained by David.

positos in domo Domini et sub manibus sacerdotum ac Leuitarum quos distribuit Dauid in domo Domini ut offerrent holocausta Domino sicut scriptum est in lege Mosi in gaudio et canticis iuxta dispositionem Dauid 23:19 constituit quo- 23:19 And he set the porters at the gates of the house of the

LORD, that none which was unclean in any thing should enter in.

que ianitores in portis domus Domini ut non ingrederetur eam inmundus in omni re 23:20 adsumpsitque centuriones 23:20 And he took the captains of hundreds, and the no-

bles, and the governors of the people, and all the people of the land, and brought down the king from the house of the LORD: and they came through the high gate into the king’s house, and set the king upon the throne of the kingdom.

et fortissimos uiros ac principes populi et omne uulgus terrae et fecerunt descendere regem de domo Domini et introire per medium portae superioris in domum regis et conlocauerunt eum in solio regali 23:21 laetatusque est omnis populus ter- 23:21 And all the people of the land rejoiced: and the

city was quiet, after that they had slain Athaliah with the sword.

rae et urbs quieuit porro Otholia interfecta est gladio

24:1 septem annorum erat Ioas cum regnare coepisset et 24:1 Joash was seven years old when he began to reign,

and he reigned forty years in Jerusalem. His mother’s name also was Zibiah of Beersheba.

quadraginta annis regnauit in Hierusalem nomen matris eius

Sebia de Bersabee 24:2 fecitque quod bonum est coram 24:2 And Joash did that which was right in the sight of the LORD all the days of Jehoiada the priest.

Biblia Sacra Vulgata

24:3 And Jehoiada took for him two wives; and he begat Domino cunctis diebus Ioiadae sacerdotis 24:3 accepit au-

sons and daughters.

tem ei Ioiadae uxores duas e quibus genuit filios et filias

24:4 And it came to pass after this, that Joash was minded 24:4 post quae placuit Ioas ut instauraret domum Domini to repair the house of the LORD.

24:5 And he gathered together the priests and the Levites, 24:5 congregauitque sacerdotes et Leuitas et dixit eis eg-

and said to them, Go out unto the cities of Judah, and gather of all Israel money to repair the house of your God from year to year, and see that ye hasten the matter. Howbeit the Levites hastened it not.

redimini ad ciuitates Iuda et colligite de uniuerso Israhel pecuniam ad sarta tecta templi Dei uestri per singulos annos festinatoque hoc facite porro Leuitae egere neglegen-

24:6 And the king called for Jehoiada the chief, and said tius 24:6 uocauitque rex Ioiadae principem et dixit ei quare

unto him, Why hast thou not required of the Levites to bring in out of Judah and out of Jerusalem the collection, according to the commandment of Moses the servant of the LORD, and of the congregation of Israel, for the tabernacle of witness?

non tibi fuit curae ut cogeres Leuitas inferre de Iuda et de Hierusalem pecuniam quae constituta est a Mose seruo Domini ut inferret eam omnis multitudo Israhel in tabernacu-

24:7 For the sons of Athaliah, that wicked woman, had lum testimonii 24:7 Otholia enim impiissima et filii eius

broken up the house of God; and also all the dedicated things of the house of the LORD did they bestow upon

Baalim. destruxerunt domum Domini et de uniuersis quae sanctifi-24:8 And at the king’s commandment they made a chest, cata fuerant templo Domini ornauerunt fanum Baalim 24:8

and set it without at the gate of the house of the LORD.

praecepit ergo rex et fecerunt arcam posueruntque eam iuxta

24:9 And they made a proclamation through Judah and portam domus Domini forinsecus 24:9 et praedicatum est

Jerusalem, to bring in to the LORD the collection that Moses the servant of God laid upon Israel in the wilder-

ness. in Iuda et Hierusalem ut deferrent singuli pretium Domino

quod constituit Moses seruus Dei super omnem Israhel in

24:10 And all the princes and all the people rejoiced, and deserto 24:10 laetatique sunt cuncti principes et omnis po-

brought in, and cast into the chest, until they had made an end.

pulus et ingressi contulerunt in arcam Domini atque mise-

24:11 Now it came to pass, that at what time the chest was runt ita ut impleretur 24:11 cumque tempus esset ut defer-

brought unto the king’s office by the hand of the Levites, and when they saw that there was much money, the king’s scribe and the high priest’s officer came and emptied the chest, and took it, and carried it to his place again. Thus they did day by day, and gathered money in abundance.

rent arcam coram rege per manus Leuitarum uidebant enim multam pecuniam ingrediebatur scriba regis et quem primus sacerdos constituerat effundebantque pecuniam quae erat in arca porro arcam reportabant ad locum suum sicque faciebant per singulos dies et congregata est infinita pecunia

24:12 And the king and Jehoiada gave it to such as did the 24:12 quam dederunt rex et Ioiada his qui praeerant operibus

work of the service of the house of the LORD, and hired masons and carpenters to repair the house of the LORD, and also such as wrought iron and brass to mend the house of the LORD.

domus Domini at illi conducebant ex ea caesores lapidum et artifices operum singulorum ut instaurarent domum Domini fabros quoque ferri et aeris ut quod cadere coeperat

24:13 So the workmen wrought, and the work was per- fulciretur 24:13 egeruntque hii qui operabantur industrie et

fected by them, and they set the house of God in his state, and strengthened it.

obducebatur parietum cicatrix per manus eorum ac suscitauerunt domum Domini in statum pristinum et firme eam

Liber II Paralipomenon

stare fecerunt 24:14 cumque conplessent omnia opera de- 24:14 And when they had finished it, they brought the rest

of the money before the king and Jehoiada, whereof were made vessels for the house of the LORD, even vessels to minister, and to offer withal, and spoons, and vessels of gold and silver. And they offered burnt offerings in the house of the LORD continually all the days of Jehoiada.

tulerunt coram rege et Ioiadae reliquam partem pecuniae de qua facta sunt uasa templi in ministerium et ad holocausta fialae quoque et cetera uasa aurea et argentea et offerebantur holocausta in domo Domini iugiter cunctis diebus Ioiadae 24:15 senuit autem Ioiadae plenus dierum et mortuus est cum 24:15 But Jehoiada waxed old, and was full of days when

he died; an hundred and thirty years old was he when he died.

centum triginta esset annorum 24:16 sepelieruntque eum in

24:16 And they buried him in the city of David among the kings, because he had done good in Israel, both toward God, and toward his house.

ciuitate Dauid cum regibus eo quod fecisset bonum cum Israhel et cum domo eius 24:17 postquam autem obiit Ioiada 24:17 Now after the death of Jehoiada came the princes

of Judah, and made obeisance to the king. Then the king hearkened unto them.

ingressi sunt principes Iuda et adorauerunt regem qui delinitus obsequiis eorum adquieuit eis 24:18 et dereliquerunt 24:18 And they left the house of the LORD God of their

fathers, and served groves and idols: and wrath came upon Judah and Jerusalem for this their trespass.

templum Domini Dei patrum suorum seruieruntque lucis et sculptilibus et facta est ira contra Iudam et Hierusalem propter hoc peccatum 24:19 mittebatque eis prophetas ut reuer- 24:19 Yet he sent prophets to them, to bring them again

unto the LORD; and they testified against them: but they would not give ear.

terentur ad Dominum quos protestantes illi audire nolebant 24:20 spiritus itaque Dei induit Zacchariam filium Ioiadae sa- 24:20 And the Spirit of God came upon Zechariah the son

of Jehoiada the priest, which stood above the people, and said unto them, Thus saith God, Why transgress ye the commandments of the LORD, that ye cannot prosper? because ye have forsaken the LORD, he hath also forsaken you.

cerdotem et stetit in conspectu populi et dixit eis haec dicit Dominus quare transgredimini praeceptum Domini quod uobis non prodeeit et dereliquistis Dominum ut derelinqueret uos 24:21 qui congregati aduersus eum miserunt lapides 24:21 And they conspired against him, and stoned him

with stones at the commandment of the king in the court of the house of the LORD.

iuxta regis imperium in atrio domus Domini 24:22 et non

24:22 Thus Joash the king remembered not the kindness which Jehoiada his father had done to him, but slew his son. And when he died, he said, The LORD look upon it, and require it.

est recordatus Ioas rex misericordiae quam fecerat Ioiadae pater illius secum sed interfecit filium eius qui cum moreretur ait uideat Dominus et requirat 24:23 cumque euolutus 24:23 And it came to pass at the end of the year, that the

host of Syria came up against him: and they came to Judah and Jerusalem, and destroyed all the princes of the people from among the people, and sent all the spoil of them unto the king of Damascus.

esset annus ascendit contra eum exercitus Syriae uenitque in Iudam et Hierusalem et interfecit cunctos principes populi atque uniuersam praedam miserunt regi Damascum 24:24 24:24 For the army of the Syrians came with a small com-

pany of men, and the LORD delivered a very great host into their hand, because they had forsaken the LORD God

et certe cum permodicus uenisset numerus Syrorum tradidit of their fathers. So they executed judgment against Joash. Dominus manibus eorum infinitam multitudinem eo quod reliquissent Dominum Deum patrum suorum in Ioas quoque ignominiosa exercuere iudicia 24:25 et abeuntes dimi- 24:25 And when they were departed from him, (for they

left him in great diseases,) his own servants conspired against him for the blood of the sons of Jehoiada the priest, and slew him on his bed, and he died: and they buried him in the city of David, but they buried him not in the sepulchres of the kings.

serunt eum in languoribus magnis surrexerunt autem contra eum serui sui in ultionem sanguinis filii Ioiadae sacerdotis et

Biblia Sacra Vulgata

occiderunt eum in lectulo suo et mortuus est sepelieruntque

24:26 And these are they that conspired against him; eum in ciuitate Dauid sed non in sepulchris regum 24:26

Zabad the son of Shimeath an Ammonitess, and Jehozabad the son of Shimrith a Moabitess.

insidiati uero sunt ei Zabath filius Semath Ammanitidis et

24:27 Now concerning his sons, and the greatness of the Iozabath filius Semarith Moabitidis 24:27 porro filii eius ac

burdens laid upon him, and the repairing of the house of God, behold, they are written in the story of the book of

the kings. And Amaziah his son reigned in his stead. summa pecuniae quae adunata fuerat sub eo et instauratio domus Dei scripta sunt diligentius in libro regum regnauitque Amasias filius eius pro eo

25:1 Amaziah was twenty and five years old when he be- 25:1 uiginti quinque annorum erat Amasias cum regnare

gan to reign, and he reigned twenty and nine years in Jerusalem. And his mother’s name was Jehoaddan of

Jerusalem. coepisset et uiginti nouem annis regnauit in Hierusalem no-25:2 And he did that which was right in the sight of the men matris eius Ioaden de Hierusalem 25:2 fecitque bo-

LORD, but not with a perfect heart.

num in conspectu Domini uerumtamen non in corde perfecto

25:3 Now it came to pass, when the kingdom was estab- 25:3 cumque roboratum sibi uideret imperium iugulauit ser-

lished to him, that he slew his servants that had killed the king his father.

uos qui occiderant regem patrem suum 25:4 sed filios eo-

25:4 But he slew not their children, but did as it is written in the law in the book of Moses, where the LORD commanded, saying, The fathers shall not die for the children, neither shall the children die for the fathers, but every man shall die for his own sin.

rum non interfecit sicut scriptum est in libro legis Mosi ubi praecepit Dominus dicens non occidentur patres pro filiis neque filii pro patribus suis sed unusquisque in suo peccato

25:5 Moreover Amaziah gathered Judah together, and morietur 25:5 congregauit igitur Amasias Iudam et consti-

made them captains over thousands, and captains over hundreds, according to the houses of their fathers, throughout all Judah and Benjamin: and he numbered them from twenty years old and above, and found them three hundred thousand choice men, able to go forth to war, that could handle spear and shield.

tuit eos per familias tribunosque et centuriones in uniuerso Iuda et Beniamin et recensuit a uiginti annis sursum inuenitque triginta milia iuuenum qui egrederentur ad pugnam et

25:6 He hired also an hundred thousand mighty men of tenerent hastam et clypeum 25:6 mercede quoque conduxit

valour out of Israel for an hundred talents of silver.

de Israhel centum milia robustorum centum talentis argenti

25:7 But there came a man of God to him, saying, O king, 25:7 uenit autem homo Dei ad illum et ait o rex ne egrediatur

let not the army of Israel go with thee; for the LORD is not with Israel, to wit, with all the children of Ephraim.

tecum exercitus Israhel non est enim Dominus cum Israhel

25:8 But if thou wilt go, do it; be strong for the battle: God et cunctis filiis Ephraim 25:8 quod si putas in robore exshall make thee fall before the enemy: for God hath power to help, and to cast down.

ercitus bella consistere superari te faciet Deus ab hostibus

25:9 And Amaziah said to the man of God, But what shall Dei quippe est et adiuuare et in fugam uertere 25:9 dixit-

we do for the hundred talents which I have given to the army of Israel? And the man of God answered, The LORD

is able to give thee much more than this. que Amasias ad hominem Dei quid ergo fiet de centum ta-

lentis quae dedi militibus Israhel et respondit ei homo Dei

25:10 Then Amaziah separated them, to wit, the army habet Dominus unde tibi dare possit multo his plura 25:10

that was come to him out of Ephraim, to go home again: wherefore their anger was greatly kindled against Judah,

and they returned home in great anger. separauit itaque Amasias exercitum qui uenerat ad eum ex

Ephraim ut reuerteretur in locum suum at illi contra Iudam

25:11 And Amaziah strengthened himself, and led forth uehementer irati reuersi sunt in regionem suam 25:11 porro

his people, and went to the valley of salt, and smote of the children of Seir ten thousand.

Liber II Paralipomenon

Amasias confidenter eduxit populum suum et abiit in ual-

lem Salinarum percussitque filios Seir decem milia 25:12 25:12 And other ten thousand left alive did the children of Judah carry away captive, and brought them unto the top of the rock, and cast them down from the top of the rock,

et alia decem milia uirorum ceperunt filii Iuda et adduxerunt that they all were broken in pieces. ad praeruptum cuiusdam petrae praecipitaueruntque eos de summo in praeceps qui uniuersi crepuerunt 25:13 at ille ex- 25:13 But the soldiers of the army which Amaziah sent

back, that they should not go with him to battle, fell upon the cities of Judah, from Samaria even unto Bethhoron,

ercitus quem remiserat Amasias ne secum iret ad proelium and smote three thousand of them, and took much spoil. diffusus est in ciuitatibus Iuda a Samaria usque Bethoron et interfectis tribus milibus diripuit praedam magnam 25:14 25:14 Now it came to pass, after that Amaziah was come

from the slaughter of the Edomites, that he brought the gods of the children of Seir, and set them up to be his gods, and bowed down himself before them, and burned incense unto them.

Amasias uero post caedem Idumeorum et adlatos deos filiorum Seir statuit illos in deos sibi et adorabat eos et illis adolebat incensum 25:15 quam ob rem iratus Dominus contra 25:15 Wherefore the anger of the LORD was kindled

against Amaziah, and he sent unto him a prophet, which said unto him, Why hast thou sought after the gods of the people, which could not deliver their own people out of thine hand?

Amasiam misit ad illum prophetam qui diceret ei cur adorasti deos qui non liberauerunt populum suum de manu tua 25:16 cumque haec ille loqueretur respondit ei num consi- 25:16 And it came to pass, as he talked with him, that the

king said unto him, Art thou made of the king’s counsel? forbear; why shouldest thou be smitten? Then the prophet forbare, and said, I know that God hath determined to destroy thee, because thou hast done this, and hast not hearkened unto my counsel.

liarius regis es quiesce ne interficiam te discedensque propheta scio inquit quod cogitauerit Dominus occidere te qui et fecisti hoc malum et insuper non adquieuisti consilio meo 25:17 igitur Amasias rex Iuda inito pessimo consilio misit 25:17 Then Amaziah king of Judah took advice, and sent

to Joash, the son of Jehoahaz, the son of Jehu, king of Israel, saying, Come, let us see one another in the face.

ad Ioas filium Ioachaz filii Hieu regem Israhel dicens ueni uideamus nos mutuo 25:18 at ille remisit nuntium dicens 25:18 And Joash king of Israel sent to Amaziah king of

Judah, saying, The thistle that was in Lebanon sent to the cedar that was in Lebanon, saying, Give thy daughter to my son to wife: and there passed by a wild beast that was in Lebanon, and trode down the thistle.

carduus qui est in Libano misit ad cedrum Libani dicens da filiam tuam filio meo uxorem et ecce bestiae quae erant in silua Libani transierunt et conculcauerunt carduum 25:19 25:19 Thou sayest, Lo, thou hast smitten the Edomites;

and thine heart lifteth thee up to boast: abide now at home; why shouldest thou meddle to thine hurt, that thou

dixisti percussi Edom et idcirco erigitur cor tuum in super- shouldest fall, even thou, and Judah with thee? biam sede in domo tua cur malum aduersum te prouocas ut cadas et tu et Iudas tecum 25:20 noluit audire Amasias eo 25:20 But Amaziah would not hear; for it came of God,

that he might deliver them into the hand of their enemies, because they sought after the gods of Edom.

quod Domini esset uoluntas ut traderetur in manibus hostium propter deos Edom 25:21 ascendit igitur Ioas rex Israhel et 25:21 So Joash the king of Israel went up; and they saw

one another in the face, both he and Amaziah king of Judah, at Bethshemesh, which belongeth to Judah.

mutuos sibi praebuere conspectus Amasias autem rex Iuda erat in Bethsames Iudae 25:22 corruitque Iudas coram Isra- 25:22 And Judah was put to the worse before Israel, and

they fled every man to his tent.

hel et fugit in tabernacula sua 25:23 porro Amasiam regem 25:23 And Joash the king of Israel took Amaziah king

of Judah, the son of Joash, the son of Jehoahaz, at Bethshemesh, and brought him to Jerusalem, and brake down the wall of Jerusalem from the gate of Ephraim to the corner gate, four hundred cubits.

Iuda filium Ioas filii Ioachaz cepit Ioas rex Israhel in Bethsames et adduxit in Hierusalem destruxitque murum eius a

Biblia Sacra Vulgata

porta Ephraim usque ad portam Anguli quadringentis cubi-

25:24 And he took all the gold and the silver, and all the tis 25:24 omne quoque aurum et argentum et uniuersa uasa

vessels that were found in the house of God with Obededom, and the treasures of the king’s house, the hostages

also, and returned to Samaria. quae reppererat in domo Dei et apud Obededom in thesauris

etiam domus regiae necnon et filios obsidum reduxit Sama-

25:25 And Amaziah the son of Joash king of Judah lived riam 25:25 uixit autem Amasias filius Ioas rex Iuda post-

after the death of Joash son of Jehoahaz king of Israel fifteen years.

quam mortuus est Ioas filius Ioachaz rex Israhel quindecim

25:26 Now the rest of the acts of Amaziah, first and last, annis 25:26 reliqua uero sermonum Amasiae priorum et no-

behold, are they not written in the book of the kings of Judah and Israel?

uissimorum scripta sunt in libro regum Iuda et Israhel 25:27

25:27 Now after the time that Amaziah did turn away from following the LORD they made a conspiracy against him in Jerusalem; and he fled to Lachish: but they sent to Lachish after him, and slew him there.

qui postquam recessit a Domino tetenderunt ei insidias in Hierusalem cumque fugisset Lachis miserunt et interfece-

25:28 And they brought him upon horses, and buried him runt eum ibi 25:28 reportantesque super equos sepelierunt

with his fathers in the city of Judah.

eum cum patribus suis in ciuitate Dauid

26:1 Then all the people of Judah took Uzziah, who was 26:1 omnis autem populus Iuda filium eius Oziam annosixteen years old, and made him king in the room of his father Amaziah.

rum sedecim constituit regem pro patre suo Amasia 26:2

26:2 He built Eloth, and restored it to Judah, after that the king slept with his fathers.

ipse aedificauit Ahilath et restituit eam dicioni Iudae post-

26:3 Sixteen years old was Uzziah when he began to reign, quam dormiuit rex cum patribus suis 26:3 sedecim annoand he reigned fifty and two years in Jerusalem. His mother’s name also was Jecoliah of Jerusalem.

rum erat Ozias cum regnare coepisset et quinquaginta duobus annis regnauit in Hierusalem nomen matris eius Hieche-

26:4 And he did that which was right in the sight of the lia de Hierusalem 26:4 fecitque quod erat rectum in oculis

LORD, according to all that his father Amaziah did.

26:5 And he sought God in the days of Zechariah, who Domini iuxta omnia quae fecerat Amasias pater eius 26:5 et

had understanding in the visions of God: and as long as he sought the LORD, God made him to prosper.

exquisiuit Deum in diebus Zacchariae intellegentis et uidentis Deum cumque requireret Dominum direxit eum in omni-

26:6 And he went forth and warred against the Philistines, bus 26:6 denique egressus est et pugnauit contra Philisthim

and brake down the wall of Gath, and the wall of Jabneh, and the wall of Ashdod, and built cities about Ashdod, and

among the Philistines. et destruxit murum Geth et murum Iabniae murumque Azoti 26:7 And God helped him against the Philistines, and aedificauit quoque oppida in Azoto et in Philisthim 26:7 et

against the Arabians that dwelt in Gurbaal, and the Mehunims.

adiuuit eum Deus contra Philisthim et contra Arabas qui ha-

26:8 And the Ammonites gave gifts to Uzziah: and his bitabant in Gurbaal et contra Ammanitas 26:8 pendebant-

name spread abroad even to the entering in of Egypt; for he strengthened himself exceedingly.

que Ammanitae munera Oziae et diuulgatum est nomen eius

26:9 Moreover Uzziah built towers in Jerusalem at the cor- usque ad introitum Aegypti propter crebras uictorias 26:9

ner gate, and at the valley gate, and at the turning of the wall, and fortified them.

aedificauitque Ozias turres in Hierusalem super portam Anguli et super portam Vallis et reliquas in eodem muri latere

26:10 Also he built towers in the desert, and digged many firmauitque eas 26:10 extruxit etiam turres in solitudine et

wells: for he had much cattle, both in the low country, and in the plains: husbandmen also, and vine dressers in the

mountains, and in Carmel: for he loved husbandry. fodit cisternas plurimas eo quod haberet multa pecora tam

Liber II Paralipomenon

in campestribus quam in heremi uastitate uineas quoque habuit et uinitores in montibus et in Carmelo erat quippe homo agriculturae deditus 26:11 fuit autem exercitus bellatorum 26:11 Moreover Uzziah had an host of fighting men, that

went out to war by bands, according to the number of their account by the hand of Jeiel the scribe and Maaseiah the ruler, under the hand of Hananiah, one of the king’s captains.

eius qui procedebant ad proelia sub manu Hiehihel scribae Maasiaeque doctoris et sub manu Ananiae qui erat de ducibus regis 26:12 omnisque numerus principum per familias 26:12 The whole number of the chief of the fathers of the

mighty men of valour were two thousand and six hundred.

uirorum fortium duum milium sescentorum 26:13 et sub eis 26:13 And under their hand was an army, three hundred

thousand and seven thousand and five hundred, that made war with mighty power, to help the king against the enemy.

uniuersus exercitus trecentorum et septem milium quingentorum qui erant apti ad bella et pro rege contra aduersarios dimicabant 26:14 praeparauit quoque eis Ozias id est 26:14 And Uzziah prepared for them throughout all the

host shields, and spears, and helmets, and habergeons, and bows, and slings to cast stones.

cuncto exercitui clypeos et hastas et galeas et loricas arcusque et fundas ad iaciendos lapides 26:15 et fecit in Hieru- 26:15 And he made in Jerusalem engines, invented by cun-

ning men, to be on the towers and upon the bulwarks, to shoot arrows and great stones withal. And his name spread far abroad; for he was marvellously helped, till he was strong.

salem diuersi generis machinas quas in turribus conlocauit et in angulis murorum ut mitterent sagittas et saxa grandia egressumque est nomen eius procul eo quod auxiliaretur ei Dominus et corroborasset illum 26:16 sed cum roboratus 26:16 But when he was strong, his heart was lifted up to

his destruction: for he transgressed against the LORD his God, and went into the temple of the LORD to burn in-

esset eleuatum est cor eius in interitum suum et neglexit Do- cense upon the altar of incense. minum Deum suum ingressusque templum Domini adolere uoluit incensum super altare thymiamatis 26:17 statimque 26:17 And Azariah the priest went in after him, and with

him fourscore priests of the LORD, that were valiant men:

ingressus post eum Azarias sacerdos et cum eo sacerdotes Domini octoginta uiri fortissimi 26:18 restiterunt regi atque 26:18 And they withstood Uzziah the king, and said unto

him, It appertaineth not unto thee, Uzziah, to burn incense unto the LORD, but to the priests the sons of Aaron, that are consecrated to burn incense: go out of the sanctuary; for thou hast trespassed; neither shall it be for thine honour from the LORD God.

dixerunt non est tui officii Ozia ut adoleas incensum Domino sed sacerdotum hoc est filiorum Aaron qui consecrati sunt ad huiuscemodi ministerium egredere de sanctuario ne contempseris quia non reputabitur tibi in gloriam hoc a Domino Deo 26:19 iratusque est Ozias et tenens in manu turibulum ut ad- 26:19 Then Uzziah was wroth, and had a censer in his

hand to burn incense: and while he was wroth with the priests, the leprosy even rose up in his forehead before the priests in the house of the LORD, from beside the incense altar.

oleret incensum minabatur sacerdotibus statimque orta est lepra in fronte eius coram sacerdotibus in domo Domini super altare thymiamatis 26:20 cumque respexisset eum Aza- 26:20 And Azariah the chief priest, and all the priests,

looked upon him, and, behold, he was leprous in his forehead, and they thrust him out from thence; yea, himself

rias pontifex et omnes reliqui sacerdotes uiderunt lepram in hasted also to go out, because the LORD had smitten him. fronte eius et festinato expulerunt eum sed et ipse perterritus adcelerauit egredi eo quod sensisset ilico plagam Domini 26:21 fuit igitur Ozias rex leprosus usque ad diem mortis suae 26:21 And Uzziah the king was a leper unto the day of his

death, and dwelt in a several house, being a leper; for he was cut off from the house of the LORD: and Jotham his son was over the king’s house, judging the people of the land.

Biblia Sacra Vulgata

et habitauit in domo separata plenus lepra ob quam et eiectus fuerat de domo Domini porro Ioatham filius eius rexit do-

26:22 Now the rest of the acts of Uzziah, first and last, did mum regis et iudicabat populum terrae 26:22 reliqua autem

Isaiah the prophet, the son of Amoz, write.

sermonum Oziae priorum et nouissimorum scripsit Esaias fi-

26:23 So Uzziah slept with his fathers, and they buried him lius Amos propheta 26:23 dormiuitque Ozias cum patribus with his fathers in the field of the burial which belonged to the kings; for they said, He is a leper: and Jotham his son

reigned in his stead. suis et sepelierunt eum in agro regalium sepulchrorum eo quod esset leprosus regnauitque Ioatham filius eius pro eo

27:1 Jotham was twenty and five years old when he began 27:1 uiginti quinque annorum erat Ioatham cum regnare

to reign, and he reigned sixteen years in Jerusalem. His mother’s name also was Jerushah, the daughter of Zadok.

coepisset et sedecim annis regnauit in Hierusalem nomen

27:2 And he did that which was right in the sight of the matris eius Hierusa filia Sadoc 27:2 fecitque quod rectum

LORD, according to all that his father Uzziah did: howbeit he entered not into the temple of the LORD. And the

people did yet corruptly. erat coram Domino iuxta omnia quae fecerat Ozias pater

suus excepto quod non est ingressus templum Domini et ad-

27:3 He built the high gate of the house of the LORD, and huc populus delinquebat 27:3 ipse aedificauit portam do-

on the wall of Ophel he built much.

mus Domini Excelsam et in muro Ophel multa construxit

27:4 Moreover he built cities in the mountains of Judah, 27:4 urbes quoque aedificauit in montibus Iuda et in saltibus

and in the forests he built castles and towers.

27:5 He fought also with the king of the Ammonites, and castella et turres 27:5 ipse pugnauit contra regem filiorum

prevailed against them. And the children of Ammon gave him the same year an hundred talents of silver, and ten thousand measures of wheat, and ten thousand of barley. So much did the children of Ammon pay unto him, both the second year, and the third.

Ammon et uicit eos dederuntque ei filii Ammon in tempore illo centum talenta argenti et decem milia choros tritici ac totidem choros hordei haec ei praebuerunt filii Ammon in

27:6 So Jotham became mighty, because he prepared his anno secundo et tertio 27:6 corroboratusque est Ioatham eo

ways before the LORD his God.

27:7 Now the rest of the acts of Jotham, and all his wars, quod direxisset uias suas coram Domino Deo suo 27:7 reli-

and his ways, lo, they are written in the book of the kings of Israel and Judah.

qua autem sermonum Ioatham et omnes pugnae eius et opera

27:8 He was five and twenty years old when he began to scripta sunt in libro regum Israhel et Iuda 27:8 uiginti quin-

reign, and reigned sixteen years in Jerusalem.

que annorum erat cum regnare coepisset et sedecim annis

27:9 And Jotham slept with his fathers, and they buried regnauit in Hierusalem 27:9 dormiuitque Ioatham cum pa-

him in the city of David: and Ahaz his son reigned in his stead.

tribus suis et sepelierunt eum in ciuitate Dauid et regnauit Achaz filius eius pro eo

28:1 Ahaz was twenty years old when he began to reign, 28:1 uiginti annorum erat Achaz cum regnare coepisset et

and he reigned sixteen years in Jerusalem: but he did not that which was right in the sight of the LORD, like David

his father: sedecim annis regnauit in Hierusalem non fecit rectum in 28:2 For he walked in the ways of the kings of Israel, and conspectu Domini sicut Dauid pater eius 28:2 sed ambula-

made also molten images for Baalim.

28:3 Moreover he burnt incense in the valley of the son of uit in uiis regum Israhel insuper et statuas fudit Baalim 28:3

Hinnom, and burnt his children in the fire, after the abominations of the heathen whom the LORD had cast out be-

fore the children of Israel. ipse est qui adoleuit incensum in ualle Benennon et lustrauit filios suos in igne iuxta ritum gentium quas interfecit Domi-

Liber II Paralipomenon

nus in aduentu filiorum Israhel 28:4 sacrificabat quoque et 28:4 He sacrificed also and burnt incense in the high

places, and on the hills, and under every green tree.

thymiama succendebat in excelsis et in collibus et sub omni ligno frondoso 28:5 tradiditque eum Dominus Deus eius in 28:5 Wherefore the LORD his God delivered him into the

hand of the king of Syria; and they smote him, and carried away a great multitude of them captives, and brought them to Damascus. And he was also delivered into the hand of the king of Israel, who smote him with a great slaughter.

manu regis Syriae qui percussit eum magnamque praedam de eius cepit imperio et adduxit in Damascum manibus quoque regis Israhel traditus est et percussus plaga grandi 28:6 28:6 For Pekah the son of Remaliah slew in Judah an

hundred and twenty thousand in one day, which were all valiant men; because they had forsaken the LORD God of

occiditque Phacee filius Romeliae de Iuda centum uiginti their fathers. milia in die uno omnes uiros bellatores eo quod reliquissent

Dominum Deum patrum suorum 28:7 eodem tempore oc- 28:7 And Zichri, a mighty man of Ephraim, slew Maaseiah the king’s son, and Azrikam the governor of the house, and Elkanah that was next to the king.

cidit Zechri uir potens ex Ephraim Masiam filium regis et Ezricam ducem domus eius Helcanam quoque secundum a rege 28:8 ceperuntque filii Israhel de fratribus suis ducenta 28:8 And the children of Israel carried away captive of

their brethren two hundred thousand, women, sons, and daughters, and took also away much spoil from them, and

milia mulierum puerorum et puellarum et infinitam praedam brought the spoil to Samaria. pertuleruntque eam in Samariam 28:9 ea tempestate erat ibi 28:9 But a prophet of the LORD was there, whose name

was Oded: and he went out before the host that came to Samaria, and said unto them, Behold, because the LORD God of your fathers was wroth with Judah, he hath delivered them into your hand, and ye have slain them in a rage that reacheth up unto heaven.

propheta Domini nomine Oded qui egressus obuiam exercitui uenientium in Samariam dixit eis ecce iratus Dominus Deus patrum uestrorum contra Iudam tradidit eos manibus uestris et occidistis illos atrociter ita ut caelum pertingeret uestra crudelitas 28:10 insuper filios Iuda et Hierusalem 28:10 And now ye purpose to keep under the children of

Judah and Jerusalem for bondmen and bondwomen unto you: but are there not with you, even with you, sins against

uultis uobis subicere in seruos et ancillas quod nequaquam the LORD your God? facto opus est peccatis enim super hoc Domino Deo uestro 28:11 sed audite consilium meum et reducite captiuos quos 28:11 Now hear me therefore, and deliver the captives

again, which ye have taken captive of your brethren: for the fierce wrath of the LORD is upon you.

adduxistis de fratribus uestris quia magnus furor Domini inminet uobis 28:12 steterunt itaque uiri de principibus filio- 28:12 Then certain of the heads of the children of

Ephraim, Azariah the son of Johanan, Berechiah the son of Meshillemoth, and Jehizkiah the son of Shallum, and Amasa the son of Hadlai, stood up against them that came from the war,

rum Ephraim Azarias filius Iohanan Barachias filius Mosollamoth Hiezechias filius Sellum et Amasa filius Adali contra eos qui ueniebant de proelio 28:13 et dixerunt eis non intro- 28:13 And said unto them, Ye shall not bring in the cap-

tives hither: for whereas we have offended against the LORD already, ye intend to add more to our sins and to our trespass: for our trespass is great, and there is fierce wrath against Israel.

ducetis huc captiuos ne peccemus Domino quare uultis adicere super peccata nostra et uetera cumulare delicta grande quippe peccatum est et ira furoris Domini inminet super Israhel 28:14 dimiseruntque uiri bellatores praedam et uniuersa 28:14 So the armed men left the captives and the spoil

before the princes and all the congregation.

quae ceperant coram principibus et omni multitudine 28:15 28:15 And the men which were expressed by name rose

up, and took the captives, and with the spoil clothed all that were naked among them, and arrayed them, and shod them, and gave them to eat and to drink, and anointed them, and carried all the feeble of them upon asses, and brought them to Jericho, the city of palm trees, to their brethren: then they returned to Samaria.

steteruntque uiri quos supra memorauimus et adprehenden-

Biblia Sacra Vulgata

tes captiuos omnesque qui nudi erant uestierunt de spoliis cumque uestissent eos et calciassent et refecissent cibo ac potu unxissent quoque propter laborem et adhibuissent eis curam quicumque ambulare non poterant et erant inbecillo corpore inposuerunt eos iumentis et adduxerunt Hierichum ciuitatem Palmarum ad fratres eorum ipsique reuersi sunt

28:16 At that time did king Ahaz send unto the kings of Samariam 28:16 tempore illo misit rex Achaz ad regem As-

Assyria to help him.

28:17 For again the Edomites had come and smitten Judah, syriorum auxilium postulans 28:17 ueneruntque Idumei et

and carried away captives.

percusserunt multos ex Iuda et ceperunt praedam magnam

28:18 The Philistines also had invaded the cities of the 28:18 Philisthim quoque diffusi sunt per urbes campestres et

low country, and of the south of Judah, and had taken Bethshemesh, and Ajalon, and Gederoth, and Shocho with the villages thereof, and Timnah with the villages thereof, Gimzo also and the villages thereof: and they dwelt there.

ad meridiem Iuda ceperuntque Bethsames et Ahilon et Gaderoth Soccho quoque et Thamnam et Gamzo cum uiculis

28:19 For the LORD brought Judah low because of Ahaz suis et habitauerunt in eis 28:19 humiliauerat enim Domi-

king of Israel; for he made Judah naked, and transgressed sore against the LORD.

nus Iudam propter Achaz regem Iuda eo quod nudasset eum

28:20 And Tilgathpilneser king of Assyria came unto him, auxilio et contemptui habuisset Dominum 28:20 adduxitand distressed him, but strengthened him not.

que contra eum Thaglathphalnasar regem Assyriorum qui

28:21 For Ahaz took away a portion out of the house of et adflixit eum et nullo resistente uastauit 28:21 igitur Ac-

the LORD, and out of the house of the king, and of the princes, and gave it unto the king of Assyria: but he helped

him not. haz spoliata domo Domini et domo regum et principum de-28:22 And in the time of his distress did he trespass yet dit regi Assyriorum munera et tamen nihil ei profuit 28:22

more against the LORD: this is that king Ahaz.

insuper et in tempore angustiae suae auxit contemptum in

28:23 For he sacrificed unto the gods of Damascus, which Dominum ipse per se rex Achaz 28:23 immolauit diis Dasmote him: and he said, Because the gods of the kings of Syria help them, therefore will I sacrifice to them, that they may help me. But they were the ruin of him, and of all Israel.

masci uictimas percussoribus suis et dixit dii regum Syriae auxiliantur eis quos ego placabo hostiis et aderunt mihi cum

28:24 And Ahaz gathered together the vessels of the house e contrario ipsi fuerint ruina eius et uniuerso Israhel 28:24

of God, and cut in pieces the vessels of the house of God, and shut up the doors of the house of the LORD, and he

made him altars in every corner of Jerusalem. direptis itaque Achaz omnibus uasis domus Dei atque con-

fractis clusit ianuas templi Dei et fecit sibi altaria in uniuersis

28:25 And in every several city of Judah he made high angulis Hierusalem 28:25 in omnibus quoque urbibus Iuda

places to burn incense unto other gods, and provoked to anger the LORD God of his fathers.

extruxit aras ad cremandum tus atque ad iracundiam prouo-

28:26 Now the rest of his acts and of all his ways, first and cauit Dominum Deum patrum suorum 28:26 reliqua autem last, behold, they are written in the book of the kings of Judah and Israel.

sermonum eius et omnium operum priorum et nouissimorum

28:27 And Ahaz slept with his fathers, and they buried scripta sunt in libro regum Iuda et Israhel 28:27 dormiuit-

him in the city, even in Jerusalem: but they brought him not into the sepulchres of the kings of Israel: and Hezekiah

his son reigned in his stead. que Achaz cum patribus suis et sepelierunt eum in ciuitate Hierusalem neque enim receperunt eum in sepulchra regum

Liber II Paralipomenon

Israhel regnauitque Ezechias filius eius pro eo

29:1 igitur Ezechias regnare coepit cum uiginti quinque 29:1 Hezekiah began to reign when he was five and

twenty years old, and he reigned nine and twenty years in Jerusalem. And his mother’s name was Abijah, the daugh-

esset annorum et uiginti nouem annis regnauit in Hierusa- ter of Zechariah. lem nomen matris eius Abia filia Zacchariae 29:2 fecitque 29:2 And he did that which was right in the sight of the

LORD, according to all that David his father had done.

quod erat placitum in conspectu Domini iuxta omnia quae fecerat Dauid pater eius 29:3 ipse anno et mense primo re- 29:3 He in the first year of his reign, in the first month,

opened the doors of the house of the LORD, and repaired them.

gni sui aperuit ualuas domus Domini et instaurauit eas 29:4

29:4 And he brought in the priests and the Levites, and gathered them together into the east street,

adduxitque sacerdotes atque Leuitas et congregauit eos in plateam orientalem 29:5 dixitque ad eos audite me Leui- 29:5 And said unto them, Hear me, ye Levites, sanctify

now yourselves, and sanctify the house of the LORD God of your fathers, and carry forth the filthiness out of the

tae et sanctificamini mundate domum Domini Dei patrum holy place. uestrorum auferte omnem inmunditiam de sanctuario 29:6 29:6 For our fathers have trespassed, and done that which

was evil in the eyes of the LORD our God, and have forsaken him, and have turned away their faces from the habi-

peccauerunt patres nostri et fecerunt malum in conspectu tation of the LORD, and turned their backs. Domini Dei nostri derelinquentes eum auerterunt facies suas a tabernaculo Domini et praebuerunt dorsum 29:7 cluserunt 29:7 Also they have shut up the doors of the porch, and

put out the lamps, and have not burned incense nor offered burnt offerings in the holy place unto the God of Israel.

ostia quae erant in porticu et extinxerunt lucernas incensumque non adoleuerunt et holocausta non obtulerunt in sanctuario Deo Israhel 29:8 concitatus est itaque furor Domini 29:8 Wherefore the wrath of the LORD was upon Judah

and Jerusalem, and he hath delivered them to trouble, to astonishment, and to hissing, as ye see with your eyes.

super Iudam et Hierusalem tradiditque eos in commotionem et in interitum et in sibilum sicut ipsi cernitis oculis uestris 29:9 en corruerunt patres nostri gladiis filii nostri et filiae 29:9 For, lo, our fathers have fallen by the sword, and our

sons and our daughters and our wives are in captivity for this.

nostrae et coniuges captiuae ductae sunt propter hoc scelus 29:10 nunc igitur placet mihi ut ineamus foedus cum 29:10 Now it is in mine heart to make a covenant with the

LORD God of Israel, that his fierce wrath may turn away from us.

Domino Deo Israhel et auertat a nobis furorem irae suae 29:11 filii mi nolite neglegere uos elegit Dominus ut stetis 29:11 My sons, be not now negligent: for the LORD hath

chosen you to stand before him, to serve him, and that ye should minister unto him, and burn incense.

coram eo et ministretis illi colatis eum et cremetis incensum 29:12 surrexerunt ergo Leuitae Maath filius Amasiae 29:12 Then the Levites arose, Mahath the son of Ama-

sai, and Joel the son of Azariah, of the sons of the Kohathites: and of the sons of Merari, Kish the son of Abdi, and Azariah the son of Jehalelel: and of the Gershonites; Joah the son of Zimmah, and Eden the son of Joah:

et Iohel filius Azariae de filiis Caath porro de filiis Merari Cis filius Abdai et Azarias filius Iallelel de filiis autem Ger-

som Ioha filius Zemma et Eden filius Ioaha 29:13 at uero 29:13 And of the sons of Elizaphan; Shimri, and Jeiel: and

of the sons of Asaph; Zechariah, and Mattaniah:

de filiis Elisaphan Samri et Iahihel de filiis quoque Asaph Zaccharias et Mathanias 29:14 necnon de filiis Heman Ia- 29:14 And of the sons of Heman; Jehiel, and Shimei: and

of the sons of Jeduthun; Shemaiah, and Uzziel.

hihel et Semei sed et de filiis Idithun Semeias et Ozihel 29:15 congregaueruntque fratres suos et sanctificati sunt et 29:15 And they gathered their brethren, and sanctified

themselves, and came, according to the commandment of the king, by the words of the LORD, to cleanse the house of the LORD.

Biblia Sacra Vulgata

ingressi iuxta mandatum regis et imperium Domini ut expia-

29:16 And the priests went into the inner part of the house rent domum Dei 29:16 sacerdotes quoque ingressi templum of the LORD, to cleanse it, and brought out all the uncleanness that they found in the temple of the LORD into the court of the house of the LORD. And the Levites took it, to carry it out abroad into the brook Kidron.

Domini ut sanctificarent illud extulerunt omnem inmunditiam quam intro reppererant in uestibulum domus Domini quam tulerunt Leuitae et asportauerunt ad torrentem Ced-

29:17 Now they began on the first day of the first month to ron foras 29:17 coeperunt autem prima die mensis primi

sanctify, and on the eighth day of the month came they to the porch of the LORD: so they sanctified the house of the LORD in eight days; and in the sixteenth day of the first month they made an end.

mundare et in die octaua eiusdem mensis ingressi sunt porticum templi Domini expiaueruntque templum diebus octo et in die sextadecima mensis eiusdem quod coeperant im-

29:18 Then they went in to Hezekiah the king, and said, pleuerunt 29:18 ingressi quoque sunt ad Ezechiam regem

We have cleansed all the house of the LORD, and the altar of burnt offering, with all the vessels thereof, and the

shewbread table, with all the vessels thereof. et dixerunt ei sanctificauimus omnem domum Domini et al-

tare holocaustoseos uasaque eius necnon et mensam propo-

29:19 Moreover all the vessels, which king Ahaz in his sitionis cum omnibus uasis suis 29:19 cunctamque templi

reign did cast away in his transgression, have we prepared and sanctified, and, behold, they are before the altar of the

LORD. supellectilem quam polluerat rex Achaz in regno suo post-

quam praeuaricatus est et ecce exposita sunt omnia coram

29:20 Then Hezekiah the king rose early, and gathered the altari Domini 29:20 consurgensque diluculo Ezechias rex

rulers of the city, and went up to the house of the LORD.

adunauit omnes principes ciuitatis et ascendit domum Do-

29:21 And they brought seven bullocks, and seven rams, mini 29:21 obtuleruntque simul tauros septem arietes sep-

and seven lambs, and seven he goats, for a sin offering for the kingdom, and for the sanctuary, and for Judah. And he commanded the priests the sons of Aaron to offer them on the altar of the LORD.

tem agnos septem et hircos septem pro peccato pro regno pro sanctuario pro Iuda dixit quoque sacerdotibus filiis Aa-

29:22 So they killed the bullocks, and the priests received ron ut offerrent super altare Domini 29:22 mactauerunt igi-

the blood, and sprinkled it on the altar: likewise, when they had killed the rams, they sprinkled the blood upon the altar: they killed also the lambs, and they sprinkled the blood upon the altar.

tur tauros et susceperunt sacerdotes sanguinem et fuderunt illud super altare mactauerunt etiam arietes et illorum sanguinem super altare fuderunt immolauerunt agnos et fude-

29:23 And they brought forth the he goats for the sin of- runt super altare sanguinem 29:23 adplicauerunt hircos pro

fering before the king and the congregation; and they laid their hands upon them:

peccato coram rege et uniuersa multitudine inposueruntque

29:24 And the priests killed them, and they made recon- manus suas super eos 29:24 et immolauerunt illos sacerdo-

ciliation with their blood upon the altar, to make an atonement for all Israel: for the king commanded that the burnt

offering and the sin offering should be made for all Israel. tes et asperserunt sanguinem eorum altari pro piaculo uni-

uersi Israhelis pro omni quippe Israhel praeceperat rex ut

29:25 And he set the Levites in the house of the LORD holocaustum fieret et pro peccato 29:25 constituit quoque

with cymbals, with psalteries, and with harps, according to the commandment of David, and of Gad the king’s seer, and Nathan the prophet: for so was the commandment of the LORD by his prophets.

Leuitas in domo Domini cum cymbalis et psalteriis et citharis secundum dispositionem Dauid et Gad uidentis regis et Nathan prophetae siquidem Domini praeceptum fuit per ma-

Liber II Paralipomenon

num prophetarum eius 29:26 steteruntque Leuitae tenentes 29:26 And the Levites stood with the instruments of

David, and the priests with the trumpets.

organa Dauid et sacerdotes tubas 29:27 et iussit Ezechias ut 29:27 And Hezekiah commanded to offer the burnt offer-

ing upon the altar. And when the burnt offering began, the song of the LORD began also with the trumpets, and with

offerrent holocaustum super altare cumque offerrentur ho- the instruments ordained by David king of Israel. locausta coeperunt laudes canere Domino et clangere tubis atque in diuersis organis quae Dauid rex Israhel reppererat concrepare 29:28 omni autem turba adorante cantores et hii 29:28 And all the congregation worshipped, and the

singers sang, and the trumpeters sounded: and all this continued until the burnt offering was finished.

qui tenebant tubas erant in officio suo donec conpleretur holocaustum 29:29 cumque finita esset oblatio incuruatus est 29:29 And when they had made an end of offering, the

king and all that were present with him bowed themselves, and worshipped.

rex et omnes qui erant cum eo et adorauerunt 29:30 praece-

29:30 Moreover Hezekiah the king and the princes commanded the Levites to sing praise unto the LORD with the words of David, and of Asaph the seer. And they sang praises with gladness, and they bowed their heads and worshipped.

pitque Ezechias et principes Leuitis ut laudarent Dominum sermonibus Dauid et Asaph uidentis qui laudauerunt eum magna laetitia et curuato genu adorauerunt 29:31 Ezechias 29:31 Then Hezekiah answered and said, Now ye have

consecrated yourselves unto the LORD, come near and bring sacrifices and thank offerings into the house of the LORD. And the congregation brought in sacrifices and thank offerings; and as many as were of a free heart burnt offerings.

autem etiam haec addidit implestis manus uestras Domino accedite et offerte uictimas et laudes in domo Domini obtulit ergo uniuersa multitudo hostias et laudes et holocausta mente deuota 29:32 porro numerus holocaustorum quae ob- 29:32 And the number of the burnt offerings, which the

congregation brought, was threescore and ten bullocks, an hundred rams, and two hundred lambs: all these were for

tulit multitudo hic fuit tauros septuaginta arietes centum ag- a burnt offering to the LORD. nos ducentos 29:33 sanctificaueruntque Domino boues ses- 29:33 And the consecrated things were six hundred oxen

and three thousand sheep.

centos et oues tria milia 29:34 sacerdotes uero pauci erant 29:34 But the priests were too few, so that they could not

flay all the burnt offerings: wherefore their brethren the Levites did help them, till the work was ended, and until the other priests had sanctified themselves: for the Levites were more upright in heart to sanctify themselves than the priests.

nec poterant sufficere ut pelles holocaustorum detraherent unde et Leuitae fratres eorum adiuuerunt eos donec impleretur opus et sanctificarentur antistites Leuitae quippe faciliori ritu sanctificantur quam sacerdotes 29:35 fuerunt igitur 29:35 And also the burnt offerings were in abundance,

with the fat of the peace offerings, and the drink offerings for every burnt offering. So the service of the house

holocausta plurima adipes pacificorum et libamina holocau- of the LORD was set in order. storum et conpletus est cultus domus Domini 29:36 laeta- 29:36 And Hezekiah rejoiced, and all the people, that God

had prepared the people: for the thing was done suddenly.

tusque est Ezechias et omnis populus eo quod ministerium Domini esset expletum de repente quippe hoc fieri placuerat

30:1 misit quoque Ezechias ad omnem Israhel et Iudam 30:1 And Hezekiah sent to all Israel and Judah, and wrote letters also to Ephraim and Manasseh, that they should come to the house of the LORD at Jerusalem, to keep the

scripsitque epistulas ad Ephraim et Manassem ut uenirent passover unto the LORD God of Israel. ad domum Domini in Hierusalem et facerent phase Domino Deo Israhel 30:2 inito ergo consilio regis et principum et 30:2 For the king had taken counsel, and his princes, and

all the congregation in Jerusalem, to keep the passover in the second month.

uniuersi coetus Hierusalem decreuerunt ut facerent phase mense secundo 30:3 non enim occurrerant facere in tem- 30:3 For they could not keep it at that time, because the

priests had not sanctified themselves sufficiently, neither had the people gathered themselves together to Jerusalem.

Biblia Sacra Vulgata

pore suo quia sacerdotes qui possent sufficere sanctificati non fuerant et populus necdum congregatus erat in Hierusa-

30:4 And the thing pleased the king and all the congrega- lem 30:4 placuitque sermo regi et omni multitudini 30:5 tion. 30:5 So they established a decree to make proclamation throughout all Israel, from Beersheba even to Dan, that they should come to keep the passover unto the LORD God of Israel at Jerusalem: for they had not done it of a long time in such sort as it was written.

et decreuerunt ut mitterent nuntios in uniuersum Israhel de Bersabee usque Dan ut uenirent et facerent phase Domino Deo Israhel in Hierusalem multi enim non fecerant sicut lege

30:6 So the posts went with the letters from the king and praescriptum est 30:6 perrexeruntque cursores cum epistu-

his princes throughout all Israel and Judah, and according to the commandment of the king, saying, Ye children of Israel, turn again unto the LORD God of Abraham, Isaac, and Israel, and he will return to the remnant of you, that are escaped out of the hand of the kings of Assyria.

lis ex regis imperio et principum eius in uniuersum Israhel et Iudam iuxta quod rex iusserat praedicantes filii Israhel reuertimini ad Dominum Deum Abraham et Isaac et Israhel et reuertetur ad reliquias quae effugerunt manum regis As-

30:7 And be not ye like your fathers, and like your syriorum 30:7 nolite fieri sicut patres uestri et fratres qui

brethren, which trespassed against the LORD God of their fathers, who therefore gave them up to desolation, as ye

see. recesserunt a Domino Deo patrum suorum et tradidit eos in 30:8 Now be ye not stiffnecked, as your fathers were, but interitum ut ipsi cernitis 30:8 nolite indurare ceruices ue-

yield yourselves unto the LORD, and enter into his sanctuary, which he hath sanctified for ever: and serve the LORD your God, that the fierceness of his wrath may turn away from you.

stras sicut patres uestri tradite manus Domino et uenite ad sanctuarium eius quod sanctificauit in aeternum seruite Domino Deo patrum uestrorum et auertetur a uobis ira furo-

30:9 For if ye turn again unto the LORD, your brethren ris eius 30:9 si enim uos reuersi fueritis ad Dominum frat-

and your children shall find compassion before them that lead them captive, so that they shall come again into this land: for the LORD your God is gracious and merciful, and will not turn away his face from you, if ye return unto him.

res uestri et filii habebunt misericordiam coram dominis suis qui illos duxere captiuos et reuertentur in terram hanc pius enim et clemens est Dominus Deus uester et non auertet fa-

30:10 So the posts passed from city to city through the ciem suam a uobis si reuersi fueritis ad eum 30:10 igitur

country of Ephraim and Manasseh even unto Zebulun: but they laughed them to scorn, and mocked them.

cursores pergebant uelociter de ciuitate in ciuitatem per terram Ephraim et Manasse usque Zabulon illis inridentibus et

30:11 Nevertheless divers of Asher and Manasseh and of subsannantibus eos 30:11 attamen quidam uiri ex Aser et

Zebulun humbled themselves, and came to Jerusalem.

Manasse et Zabulon adquiescentes consilio uenerunt Hieru-

30:12 Also in Judah the hand of God was to give them salem 30:12 in Iuda uero facta est manus Domini ut daret

one heart to do the commandment of the king and of the princes, by the word of the LORD.

eis cor unum et facerent iuxta praeceptum regis et princi-

30:13 And there assembled at Jerusalem much people to pum uerbum Domini 30:13 congregatique sunt in Hierukeep the feast of unleavened bread in the second month, a very great congregation.

salem populi multi ut facerent sollemnitatem azymorum in

30:14 And they arose and took away the altars that were mense secundo 30:14 et surgentes destruxerunt altaria quae

in Jerusalem, and all the altars for incense took they away, and cast them into the brook Kidron.

erant in Hierusalem atque uniuersa in quibus idolis adolebatur incensum subuertentes proiecerunt in torrentem Cedron

Liber II Paralipomenon

30:15 immolauerunt autem phase quartadecima die mensis 30:15 Then they killed the passover on the fourteenth day

of the second month: and the priests and the Levites were ashamed, and sanctified themselves, and brought in the

secundi sacerdotes quoque atque Leuitae tandem sanctifi- burnt offerings into the house of the LORD. cati obtulerunt holocausta in domo Domini 30:16 steterunt- 30:16 And they stood in their place after their manner, ac-

cording to the law of Moses the man of God: the priests sprinkled the blood, which they received of the hand of the

que in ordine suo iuxta dispositionem et legem Mosi ho- Levites. minis Dei sacerdotes uero suscipiebant effundendum san-

guinem de manibus Leuitarum 30:17 eo quod multa turba 30:17 For there were many in the congregation that were not sanctified: therefore the Levites had the charge of the killing of the passovers for every one that was not clean,

sanctificata non esset et idcirco Leuitae immolarent phase to sanctify them unto the LORD. his qui non occurrerant sanctificari Domino 30:18 magna 30:18 For a multitude of the people, even many of

Ephraim, and Manasseh, Issachar, and Zebulun, had not cleansed themselves, yet did they eat the passover otherwise than it was written. But Hezekiah prayed for them, saying, The good LORD pardon every one

etiam pars populi de Ephraim et Manasse et Isachar et Zabulon quae sanctificata non fuerat comedit phase non iuxta quod scriptum est et orauit pro eis Ezechias dicens Dominus bonus propitiabitur 30:19 cunctis qui in toto corde requi- 30:19 That prepareth his heart to seek God, the LORD

God of his fathers, though he be not cleansed according to the purification of the sanctuary.

runt Dominum Deum patrum suorum et non inputabit eis quod minus sanctificati sunt 30:20 quem exaudiuit Domi- 30:20 And the LORD hearkened to Hezekiah, and healed

the people.

nus et placatus est populo 30:21 feceruntque filii Israhel qui 30:21 And the children of Israel that were present at

Jerusalem kept the feast of unleavened bread seven days with great gladness: and the Levites and the priests praised the LORD day by day, singing with loud instruments unto the LORD.

inuenti sunt in Hierusalem sollemnitatem azymorum septem diebus in laetitia magna laudantes Dominum per singulos dies Leuitae quoque et sacerdotes per organa quae suo officio congruebant 30:22 et locutus est Ezechias ad cor om- 30:22 And Hezekiah spake comfortably unto all the

Levites that taught the good knowledge of the LORD: and they did eat throughout the feast seven days, offering peace offerings, and making confession to the LORD God of their fathers.

nium Leuitarum qui habebant intellegentiam bonam super Domino et comederunt septem diebus sollemnitatis immolantes uictimas pacificorum et laudantes Dominum Deum patrum suorum 30:23 placuitque uniuersae multitudini ut 30:23 And the whole assembly took counsel to keep other

seven days: and they kept other seven days with gladness.

celebrarent etiam alios dies septem quod et fecerunt cum ingenti gaudio 30:24 Ezechias enim rex Iuda praebuerat mul- 30:24 For Hezekiah king of Judah did give to the congre-

gation a thousand bullocks and seven thousand sheep; and the princes gave to the congregation a thousand bullocks and ten thousand sheep: and a great number of priests sanctified themselves.

titudini mille tauros et septem milia ouium principes uero dederant populo tauros mille et oues decem milia sanctificata ergo est sacerdotum plurima multitudo 30:25 et hilari- 30:25 And all the congregation of Judah, with the priests

and the Levites, and all the congregation that came out of Israel, and the strangers that came out of the land of Israel,

tate perfusa omnis turba Iuda tam sacerdotum et Leuitarum and that dwelt in Judah, rejoiced. quam uniuersae frequentiae quae uenerat ex Israhel proselytorum quoque de terra Israhel et habitantium in Iuda 30:26 30:26 So there was great joy in Jerusalem: for since the

time of Solomon the son of David king of Israel there was not the like in Jerusalem.

factaque est grandis celebritas in Hierusalem qualis a diebus Salomonis filii Dauid regis Israhel in ea urbe non fuerat

Biblia Sacra Vulgata

30:27 Then the priests the Levites arose and blessed the 30:27 surrexerunt autem sacerdotes atque Leuitae benedicen-

people: and their voice was heard, and their prayer came up to his holy dwelling place, even unto heaven.

tes populo et exaudita est uox eorum peruenitque oratio in habitaculum sanctum caeli

31:1 Now when all this was finished, all Israel that were 31:1 cumque haec fuissent rite celebrata egressus est om-

present went out to the cities of Judah, and brake the images in pieces, and cut down the groves, and threw down the high places and the altars out of all Judah and Benjamin, in Ephraim also and Manasseh, until they had utterly destroyed them all. Then all the children of Israel returned, every man to his possession, into their own cities.

nis Israhel qui inuentus fuerat in urbibus Iuda et fregerunt simulacra succideruntque lucos demoliti sunt excelsa et altaria destruxerunt non solum de uniuerso Iuda et Beniamin sed de Ephraim quoque et Manasse donec penitus euerterent reuersique sunt omnes filii Israhel in possessiones et ciui-

31:2 And Hezekiah appointed the courses of the priests tates suas 31:2 Ezechias uero constituit turmas sacerdota-

and the Levites after their courses, every man according to his service, the priests and Levites for burnt offerings and for peace offerings, to minister, and to give thanks, and to praise in the gates of the tents of the LORD.

les et leuiticas per diuisiones suas unumquemque in officio proprio tam sacerdotum uidelicet quam Leuitarum ad holocausta et pacifica ut ministrarent et confiterentur canerent-

31:3 He appointed also the king’s portion of his substance que in portis castrorum Domini 31:3 pars autem regis erat for the burnt offerings, to wit, for the morning and evening burnt offerings, and the burnt offerings for the sabbaths, and for the new moons, and for the set feasts, as it is written in the law of the LORD.

ut de propria eius substantia offerretur holocaustum mane semper et uespere sabbatis quoque et kalendis et sollemnita-

31:4 Moreover he commanded the people that dwelt in tibus ceteris sicut scriptum est in lege Mosi 31:4 praecepit

Jerusalem to give the portion of the priests and the Levites, that they might be encouraged in the law of the LORD.

etiam populo habitantium Hierusalem ut darent partes sacer-

31:5 And as soon as the commandment came abroad, the dotibus et Leuitis et possent uacare legi Domini 31:5 quod

children of Israel brought in abundance the firstfruits of corn, wine, and oil, and honey, and of all the increase of the field; and the tithe of all things brought they in abundantly.

cum percrebruisset in auribus multitudinis plurimas obtulere primitias filii Israhel frumenti uini et olei mellis quoque et

31:6 And concerning the children of Israel and Judah, that omnium quae gignit humus decimas obtulerunt 31:6 sed et dwelt in the cities of Judah, they also brought in the tithe of oxen and sheep, and the tithe of holy things which were consecrated unto the LORD their God, and laid them by heaps.

filii Israhel et Iuda qui habitabant in urbibus Iuda obtulerunt decimas boum et ouium decimasque sanctorum quae uouerant Domino Deo suo atque uniuersa portantes fecerunt

31:7 In the third month they began to lay the foundation aceruos plurimos 31:7 mense tertio coeperunt aceruorum

of the heaps, and finished them in the seventh month.

31:8 And when Hezekiah and the princes came and saw iacere fundamenta et mense septimo conpleuerunt eos 31:8

the heaps, they blessed the LORD, and his people Israel.

cumque ingressi fuissent Ezechias et principes eius uiderunt

31:9 Then Hezekiah questioned with the priests and the aceruos et benedixerunt Domino ac populo Israhel 31:9 in-

Levites concerning the heaps.

terrogauitque Ezechias sacerdotes et Leuitas cur ita iacerent

31:10 And Azariah the chief priest of the house of Zadok acerui 31:10 respondit illi Azarias sacerdos primus de stirpe

answered him, and said, Since the people began to bring the offerings into the house of the LORD, we have had enough to eat, and have left plenty: for the LORD hath blessed his people; and that which is left is this great store.

Sadoc dicens ex quo coeperunt offerri primitiae in domo Domini comedimus et saturati sumus remanseruntque plurima

Liber II Paralipomenon

eo quod benedixerit Dominus populo suo reliquiarum autem copia est ista quam cernis 31:11 praecepit igitur Ezechias 31:11 Then Hezekiah commanded to prepare chambers in

the house of the LORD; and they prepared them,

ut praepararent horrea in domo Domini quod cum fecissent 31:12 intulerunt tam primitias quam decimas et quaecumque 31:12 And brought in the offerings and the tithes and

the dedicated things faithfully: over which Cononiah the Levite was ruler, and Shimei his brother was the next.

uouerant fideliter fuit autem praefectus eorum Chonenias Leuita et Semei frater eius secundus 31:13 post quem Ieihel 31:13 And Jehiel, and Azaziah, and Nahath, and Asahel,

and Jerimoth, and Jozabad, and Eliel, and Ismachiah, and Mahath, and Benaiah, were overseers under the hand of Cononiah and Shimei his brother, at the commandment of Hezekiah the king, and Azariah the ruler of the house of God.

et Azazias et Naath et Asahel et Ierimoth Iozabath quoque et Helihel et Iesmachias et Maath et Banaias praepositi sub manibus Choneniae et Semei fratris eius ex imperio Ezechiae regis et Azariae pontificis domus Domini ad quos omnia pertinebant 31:14 Core uero filius Iemna Leuites et ianitor 31:14 And Kore the son of Imnah the Levite, the porter

toward the east, was over the freewill offerings of God, to distribute the oblations of the LORD, and the most holy

orientalis portae praepositus erat his quae sponte offereban- things. tur Domino primitiisque et consecratis in sancta sanctorum 31:15 et sub cura eius Eden et Meniamin Hiesue et Sameias 31:15 And next him were Eden, and Miniamin, and

Jeshua, and Shemaiah, Amariah, and Shecaniah, in the cities of the priests, in their set office, to give to their

Amarias quoque et Sechenias in ciuitatibus sacerdotum ut brethren by courses, as well to the great as to the small: fideliter distribuerent fratribus suis partes minoribus atque maioribus 31:16 exceptis maribus ab annis tribus et supra 31:16 Beside their genealogy of males, from three years

old and upward, even unto every one that entereth into the house of the LORD, his daily portion for their service in

cunctis qui ingrediebantur templum Domini et quicquid per their charges according to their courses; dies singulos conducebat in ministerio atque obseruationibus iuxta diuisiones suas 31:17 sacerdotibus per familias et 31:17 Both to the genealogy of the priests by the house

of their fathers, and the Levites from twenty years old and upward, in their charges by their courses;

Leuitis a uicesimo anno et supra per ordines et turmas suas 31:18 uniuersaeque multitudini tam uxoribus quam liberis eo- 31:18 And to the genealogy of all their little ones, their

wives, and their sons, and their daughters, through all the congregation: for in their set office they sanctified them-

rum utriusque sexus fideliter cibi de his quae sanctificata selves in holiness: fuerant praebebantur 31:19 sed et filiorum Aaron per agros 31:19 Also of the sons of Aaron the priests, which were

in the fields of the suburbs of their cities, in every several city, the men that were expressed by name, to give portions to all the males among the priests, and to all that were reckoned by genealogies among the Levites.

et suburbana urbium singularum dispositi erant uiri qui partes distribuerent uniuerso sexui masculino de sacerdotibus et Leuitis 31:20 fecit ergo Ezechias uniuersa quae diximus 31:20 And thus did Hezekiah throughout all Judah, and

wrought that which was good and right and truth before the LORD his God.

in omni Iuda operatusque est bonum et rectum et uerum coram Domino Deo suo 31:21 in uniuersa cultura ministerii 31:21 And in every work that he began in the service of the

house of God, and in the law, and in the commandments, to seek his God, he did it with all his heart, and prospered.

domus Domini iuxta legem et caerimonias uolens requirere Deum suum in toto corde suo fecitque et prosperatus est

32:1 post quae et huiuscemodi ueritatem uenit Sennache- 32:1 After these things, and the establishment thereof,

Sennacherib king of Assyria came, and entered into Judah, and encamped against the fenced cities, and thought

rib rex Assyriorum et ingressus Iudam obsedit ciuitates mu- to win them for himself.

Biblia Sacra Vulgata

32:2 And when Hezekiah saw that Sennacherib was come, nitas uolens eas capere 32:2 quod cum uidisset Ezechias

and that he was purposed to fight against Jerusalem,

uenisse scilicet Sennacherib et totum belli impetum uerti

32:3 He took counsel with his princes and his mighty men contra Hierusalem 32:3 inito cum principibus consilio ui-

to stop the waters of the fountains which were without the city: and they did help him.

risque fortissimis ut obturarent capita fontium quae erant ex-

32:4 So there was gathered much people together, who tra urbem et hoc omnium decernente sententia 32:4 congre-

stopped all the fountains, and the brook that ran through the midst of the land, saying, Why should the kings of

Assyria come, and find much water? gauit plurimam multitudinem et obturauerunt cunctos fon-

tes et riuum qui fluebat in medio terrae dicentes ne ueniant

32:5 Also he strengthened himself, and built up all the wall reges Assyriorum et inueniant aquarum abundantiam 32:5 that was broken, and raised it up to the towers, and another wall without, and repaired Millo in the city of David, and

made darts and shields in abundance. aedificauit quoque agens industrie omnem murum qui fuerat dissipatus et extruxit turres desuper et forinsecus alterum murum instaurauitque Mello in ciuitate Dauid et fecit uni-

32:6 And he set captains of war over the people, and gath- uersi generis armaturam et clypeos 32:6 constituitque prin-

ered them together to him in the street of the gate of the city, and spake comfortably to them, saying,

cipes bellatorum in exercitu et conuocauit uniuersos in pla-

32:7 Be strong and courageous, be not afraid nor dismayed tea portae ciuitatis ac locutus est ad cor eorum dicens 32:7

for the king of Assyria, nor for all the multitude that is with him: for there be more with us than with him:

uiriliter agite et confortamini nolite timere nec paueatis regem Assyriorum et uniuersam multitudinem quae est cum

32:8 With him is an arm of flesh; but with us is the LORD eo multo enim plures nobiscum sunt quam cum illo 32:8

our God to help us, and to fight our battles. And the people rested themselves upon the words of Hezekiah king of

Judah. cum illo est brachium carneum nobiscum Dominus Deus noster qui auxiliator est noster pugnatque pro nobis confortatusque est populus huiuscemodi uerbis Ezechiae regis Iuda

32:9 After this did Sennacherib king of Assyria send his 32:9 quae postquam gesta sunt misit Sennacherib rex Assy-

servants to Jerusalem, (but he himself laid siege against Lachish, and all his power with him,) unto Hezekiah king of Judah, and unto all Judah that were at Jerusalem, saying,

riorum seruos suos Hierusalem ipse enim cum uniuerso exercitu obsidebat Lachis ad Ezechiam regem Iuda et ad om-

32:10 Thus saith Sennacherib king of Assyria, Whereon nem populum qui erat in urbe dicens 32:10 haec dicit Sen-

do ye trust, that ye abide in the siege in Jerusalem?

nacherib rex Assyriorum in quo habentes fiduciam sedetis

32:11 Doth not Hezekiah persuade you to give over your- obsessi in Hierusalem 32:11 num Ezechias decipit uos ut

selves to die by famine and by thirst, saying, The LORD our God shall deliver us out of the hand of the king of

Assyria? tradat morti in fame et siti adfirmans quod Dominus Deus 32:12 Hath not the same Hezekiah taken away his uester liberet uos de manu regis Assyriorum 32:12 num-

high places and his altars, and commanded Judah and Jerusalem, saying, Ye shall worship before one altar, and

burn incense upon it? quid non iste est Ezechias qui destruxit excelsa illius et alta-

ria et praecepit Iudae et Hierusalem dicens coram altari uno

32:13 Know ye not what I and my fathers have done unto adorabitis et in ipso conburetis incensum 32:13 an ignora-

all the people of other lands? were the gods of the nations of those lands any ways able to deliver their lands out of

mine hand? tis quae ego fecerim et patres mei cunctis terrarum populis numquid praeualuerunt dii gentium omniumque terrarum li-

Liber II Paralipomenon

berare regionem suam de manu mea 32:14 quis est de uni- 32:14 Who was there among all the gods of those nations

that my fathers utterly destroyed, that could deliver his people out of mine hand, that your God should be able

uersis diis gentium quas uastauerunt patres mei qui potuerit to deliver you out of mine hand? eruere populum suum de manu mea ut possit etiam Deus

uester eruere uos de hac manu 32:15 non uos ergo deci- 32:15 Now therefore let not Hezekiah deceive you, nor persuade you on this manner, neither yet believe him: for no god of any nation or kingdom was able to deliver his people out of mine hand, and out of the hand of my fathers: how much less shall your God deliver you out of mine hand?

piat Ezechias nec uana persuasione deludat neque credatis ei si enim nullus potuit deus cunctarum gentium atque regnorum liberare populum suum de manu mea et de manu patrum meorum consequenter nec Deus uester poterit eruere uos de hac manu 32:16 sed et alia multa locuti sunt serui 32:16 And his servants spake yet more against the LORD

God, and against his servant Hezekiah.

eius contra Dominum Deum et contra Ezechiam seruum eius 32:17 epistulas quoque scripsit plenas blasphemiae in Domi- 32:17 He wrote also letters to rail on the LORD God of

Israel, and to speak against him, saying, As the gods of the nations of other lands have not delivered their people out of mine hand, so shall not the God of Hezekiah deliver his people out of mine hand.

num Deum Israhel et locutus est aduersus eum sicut dii gentium ceterarum non potuerunt liberare populos suos de manu mea sic et Deus Ezechiae eruere non poterit populum suum de manu ista 32:18 insuper et clamore magno lingua iuda- 32:18 Then they cried with a loud voice in the Jews’

speech unto the people of Jerusalem that were on the wall, to affright them, and to trouble them; that they might take

ica contra populum qui sedebat in muris Hierusalem perso- the city. nabat ut terreret eos et caperet ciuitatem 32:19 locutusque 32:19 And they spake against the God of Jerusalem, as

against the gods of the people of the earth, which were the work of the hands of man.

est contra Deum Hierusalem sicut aduersum deos populorum terrae opera manuum hominum 32:20 orauerunt igitur 32:20 And for this cause Hezekiah the king, and the

prophet Isaiah the son of Amoz, prayed and cried to heaven.

Ezechias rex et Esaias filius Amos prophetes aduersum hanc blasphemiam ac uociferati sunt usque in caelum 32:21 et 32:21 And the LORD sent an angel, which cut off all the

mighty men of valour, and the leaders and captains in the camp of the king of Assyria. So he returned with shame of face to his own land. And when he was come into the house of his god, they that came forth of his own bowels slew him there with the sword.

misit Dominus angelum qui percussit omnem uirum robustum et bellatorem et principem exercitus regis Assyriorum reuersusque est cum ignominia in terram suam cumque ingressus esset domum dei sui filii qui egressi fuerant de utero eius interfecerunt eum gladio 32:22 saluauitque Dominus 32:22 Thus the LORD saved Hezekiah and the inhabitants

of Jerusalem from the hand of Sennacherib the king of Assyria, and from the hand of all other, and guided them

Ezechiam et habitatores Hierusalem de manu Sennacherib on every side. regis Assyriorum et de manu omnium et praestitit ei quietem per circuitum 32:23 multi etiam deferebant hostias et 32:23 And many brought gifts unto the LORD to

Jerusalem, and presents to Hezekiah king of Judah: so that he was magnified in the sight of all nations from thence-

sacrificia Domino Hierusalem et munera Ezechiae regi Iuda forth. qui exaltatus est post haec coram cunctis gentibus 32:24 in 32:24 In those days Hezekiah was sick to the death, and

prayed unto the LORD: and he spake unto him, and he gave him a sign.

diebus illis aegrotauit Ezechias usque ad mortem et orauit

Dominum exaudiuitque eum et dedit ei signum 32:25 sed 32:25 But Hezekiah rendered not again according to the benefit done unto him; for his heart was lifted up: therefore there was wrath upon him, and upon Judah and Jerusalem.

Biblia Sacra Vulgata

non iuxta beneficia quae acceperat retribuit quia eleuatum est cor eius et facta est contra eum ira et contra Iudam ac

32:26 Notwithstanding Hezekiah humbled himself for the Hierusalem 32:26 humiliatusque est postea eo quod exalta-

pride of his heart, both he and the inhabitants of Jerusalem, so that the wrath of the LORD came not upon them in the

days of Hezekiah. tum fuisset cor eius tam ipse quam habitatores Hierusalem

et idcirco non uenit super eos ira Domini in diebus Ezechiae

32:27 And Hezekiah had exceeding much riches and hon- 32:27 fuit autem Ezechias diues et inclitus ualde et thesauros

our: and he made himself treasuries for silver, and for gold, and for precious stones, and for spices, and for

shields, and for all manner of pleasant jewels; sibi plurimos congregauit argenti auri et lapidis pretiosi aro-

matum et armorum uniuersi generis et uasorum magni pretii

32:28 Storehouses also for the increase of corn, and wine, 32:28 apothecas quoque frumenti uini et olei et praesepia om-

and oil; and stalls for all manner of beasts, and cotes for flocks.

nium iumentorum caulasque pecoribus 32:29 et urbes exae-

32:29 Moreover he provided him cities, and possessions of flocks and herds in abundance: for God had given him substance very much.

dificauit habebat quippe greges ouium et armentorum innumerabiles eo quod dedisset ei Dominus substantiam multam

32:30 This same Hezekiah also stopped the upper water- nimis 32:30 ipse est Ezechias qui obturauit superiorem fon-

course of Gihon, and brought it straight down to the west side of the city of David. And Hezekiah prospered in all

his works. tem aquarum Gion et auertit eas subter ad occidentem urbis

Dauid in omnibus operibus suis fecit prospere quae uoluit

32:31 Howbeit in the business of the ambassadors of the 32:31 attamen in legatione principum Babylonis qui missi

princes of Babylon, who sent unto him to enquire of the wonder that was done in the land, God left him, to try him,

that he might know all that was in his heart. fuerant ad eum ut interrogarent de portento quod acciderat

super terram dereliquit eum Deus ut temptaretur et nota fie-

32:32 Now the rest of the acts of Hezekiah, and his good- rent omnia quae erant in corde eius 32:32 reliqua autem ser-

ness, behold, they are written in the vision of Isaiah the prophet, the son of Amoz, and in the book of the kings of

Judah and Israel. monum Ezechiae et misericordiarum eius scripta sunt in ui-

sione Esaiae filii Amos prophetae et in libro regum Iuda et

32:33 And Hezekiah slept with his fathers, and they buried Israhel 32:33 dormiuitque Ezechias cum patribus suis et se-

him in the chiefest of the sepulchres of the sons of David: and all Judah and the inhabitants of Jerusalem did him honour at his death. And Manasseh his son reigned in his stead.

pelierunt eum supra sepulchra filiorum Dauid et celebrauit eius exequias uniuersus Iuda et omnes habitatores Hierusalem regnauitque Manasses filius eius pro eo

33:1 Manasseh was twelve years old when he began to 33:1 duodecim annorum erat Manasses cum regnare coe-

reign, and he reigned fifty and five years in Jerusalem:

pisset et quinquaginta quinque annis regnauit in Hierusalem

33:2 But did that which was evil in the sight of the LORD, 33:2 fecit autem malum coram Domino iuxta abominationes

like unto the abominations of the heathen, whom the LORD had cast out before the children of Israel.

gentium quas subuertit Dominus coram filiis Israhel 33:3

33:3 For he built again the high places which Hezekiah his father had broken down, and he reared up altars for Baalim, and made groves, and worshipped all the host of heaven, and served them.

et conuersus instaurauit excelsa quae demolitus fuerat Ezechias pater eius construxitque aras Baalim et fecit lucos et

33:4 Also he built altars in the house of the LORD, adorauit omnem militiam caeli et coluit eam 33:4 aedifi-

whereof the LORD had said, In Jerusalem shall my name be for ever.

cauit quoque altaria in domo Domini de qua dixerat Domi-

Liber II Paralipomenon

nus in Hierusalem erit nomen meum in aeternum 33:5 ae- 33:5 And he built altars for all the host of heaven in the

two courts of the house of the LORD.

dificauit autem ea cuncto exercitui caeli in duobus atriis domus Domini 33:6 transireque fecit filios suos per ignem in 33:6 And he caused his children to pass through the fire in

the valley of the son of Hinnom: also he observed times, and used enchantments, and used witchcraft, and dealt with a familiar spirit, and with wizards: he wrought much evil in the sight of the LORD, to provoke him to anger.

ualle Benennon obseruabat somnia sectabatur auguria maleficis artibus inseruiebat habebat secum magos et incantatores multaque mala operatus est coram Domino ut inritaret eum 33:7 sculptile quoque et conflatile signum posuit 33:7 And he set a carved image, the idol which he had

made, in the house of God, of which God had said to David and to Solomon his son, In this house, and in Jerusalem, which I have chosen before all the tribes of Israel, will I put my name for ever:

in domo Domini de qua locutus est Dominus ad Dauid et ad Salomonem filium eius dicens in domo hac et in Hierusalem quam elegi de cunctis tribubus Israhel ponam nomen

meum in sempiternum 33:8 et mouere non faciam pedem 33:8 Neither will I any more remove the foot of Israel from out of the land which I have appointed for your fathers; so that they will take heed to do all that I have commanded them, according to the whole law and the statutes and the ordinances by the hand of Moses.

Israhel de terra quam tradidi patribus eorum ita dumtaxat si custodierint facere quae praecepi eis cunctamque legem et caerimonias atque iudicia per manum Mosi 33:9 igitur Ma- 33:9 So Manasseh made Judah and the inhabitants of

Jerusalem to err, and to do worse than the heathen, whom the LORD had destroyed before the children of Israel.

nasses seduxit Iudam et habitatores Hierusalem ut facerent malum super omnes gentes quas subuerterat Dominus a facie filiorum Israhel 33:10 locutusque est Dominus ad eum et ad 33:10 And the LORD spake to Manasseh, and to his peo-

ple: but they would not hearken.

populum illius et adtendere noluerunt 33:11 idcirco supe- 33:11 Wherefore the LORD brought upon them the cap-

tains of the host of the king of Assyria, which took Manasseh among the thorns, and bound him with fetters, and

rinduxit eis principes exercitus regis Assyriorum ceperunt- carried him to Babylon. que Manassen et uinctum catenis atque conpedibus duxerunt Babylonem 33:12 qui postquam coangustatus est ora- 33:12 And when he was in affliction, he besought the

LORD his God, and humbled himself greatly before the God of his fathers,

uit Dominum Deum suum et egit paenitentiam ualde coram Deo patrum suorum 33:13 deprecatusque est eum et obsec- 33:13 And prayed unto him: and he was intreated of

him, and heard his supplication, and brought him again to Jerusalem into his kingdom. Then Manasseh knew that

rauit intente et exaudiuit orationem eius reduxitque eum Hie- the LORD he was God. rusalem in regnum suum et cognouit Manasses quod Dominus ipse esset Deus 33:14 post haec aedificauit murum 33:14 Now after this he built a wall without the city of

David, on the west side of Gihon, in the valley, even to the entering in at the fish gate, and compassed about Ophel, and raised it up a very great height, and put captains of war in all the fenced cities of Judah.

extra ciuitatem Dauid ad occidentem Gion in conualle ab introitu portae Piscium per circuitum usque ad Ophel et exaltauit illum uehementer constituitque principes exercitus in cunctis ciuitatibus Iuda munitis 33:15 et abstulit deos alie- 33:15 And he took away the strange gods, and the idol

out of the house of the LORD, and all the altars that he had built in the mount of the house of the LORD, and in

nos et simulacrum de domo Domini aras quoque quas fecerat Jerusalem, and cast them out of the city. in monte domus Domini et in Hierusalem et proiecit omnia extra urbem 33:16 porro instaurauit altare Domini et immo- 33:16 And he repaired the altar of the LORD, and sac-

rificed thereon peace offerings and thank offerings, and commanded Judah to serve the LORD God of Israel.

Biblia Sacra Vulgata

lauit super illud uictimas et pacifica et laudem praecepitque

33:17 Nevertheless the people did sacrifice still in the high Iudae ut seruiret Domino Deo Israhel 33:17 attamen adhuc places, yet unto the LORD their God only. 33:18 Now the rest of the acts of Manasseh, and his prayer populus immolabat in excelsis Domino Deo suo 33:18 reunto his God, and the words of the seers that spake to him in the name of the LORD God of Israel, behold, they are

written in the book of the kings of Israel. liqua autem gestorum Manasse et obsecratio eius ad Deum suum uerba quoque uidentium qui loquebantur ad eum in nomine Domini Dei Israhel continentur in sermonibus regum

33:19 His prayer also, and how God was intreated of him, Israhel 33:19 oratio quoque eius et exauditio et cuncta pec-

and all his sins, and his trespass, and the places wherein he built high places, and set up groves and graven images, before he was humbled: behold, they are written among the sayings of the seers.

cata atque contemptus loca etiam in quibus aedificauit excelsa et fecit lucos et statuas antequam ageret paenitentiam

33:20 So Manasseh slept with his fathers, and they buried scripta sunt in sermonibus Ozai 33:20 dormiuit ergo Mahim in his own house: and Amon his son reigned in his stead.

nasses cum patribus suis et sepelierunt eum in domo sua reg-

33:21 Amon was two and twenty years old when he began nauitque pro eo filius eius Amon 33:21 uiginti duo annorum

to reign, and reigned two years in Jerusalem.

erat Amon cum regnare coepisset et duobus annis regnauit

33:22 But he did that which was evil in the sight of the in Hierusalem 33:22 fecitque malum in conspectu Domini

LORD, as did Manasseh his father: for Amon sacrificed unto all the carved images which Manasseh his father had

made, and served them; sicut fecerat Manasses pater eius et cunctis idolis quae Ma-33:23 And humbled not himself before the LORD, as nasses fuerat fabricatus immolauit atque seruiuit 33:23 et

Manasseh his father had humbled himself; but Amon trespassed more and more.

non est reueritus faciem Domini sicut reueritus est Manasses

33:24 And his servants conspired against him, and slew pater eius et multo maiora deliquit 33:24 cumque coniuras-

him in his own house.

sent aduersus eum serui sui interfecerunt eum in domo sua

33:25 But the people of the land slew all them that had 33:25 porro reliqua populi multitudo caesis his qui Amon per-

conspired against king Amon; and the people of the land made Josiah his son king in his stead.

cusserant constituit regem Iosiam filium eius pro eo

34:1 Josiah was eight years old when he began to reign, 34:1 octo annorum erat Iosias cum regnare coepisset et

and he reigned in Jerusalem one and thirty years.

34:2 And he did that which was right in the sight of the triginta et uno annis regnauit in Hierusalem 34:2 fecitque

LORD, and walked in the ways of David his father, and declined neither to the right hand, nor to the left.

quod erat rectum in conspectu Domini et ambulauit in uiis Dauid patris sui non declinauit neque ad dexteram neque ad

34:3 For in the eighth year of his reign, while he was yet sinistram 34:3 octauo autem anno regni sui cum adhuc es-

young, he began to seek after the God of David his father: and in the twelfth year he began to purge Judah and Jerusalem from the high places, and the groves, and the carved images, and the molten images.

set puer coepit quaerere Deum patris sui Dauid et duodecimo anno postquam coeperat mundauit Iudam et Hierusa-

34:4 And they brake down the altars of Baalim in his pres- lem ab excelsis et lucis simulacrisque et sculptilibus 34:4

ence; and the images, that were on high above them, he cut down; and the groves, and the carved images, and the molten images, he brake in pieces, and made dust of them, and strowed it upon the graves of them that had sacrificed unto them.

destruxeruntque coram eo aras Baalim et simulacra quae superposita fuerant demoliti sunt lucos etiam et sculptilia succidit atque comminuit et super tumulos eorum qui eis im-

34:5 And he burnt the bones of the priests upon their altars, molare consueuerant fragmenta dispersit 34:5 ossa praeand cleansed Judah and Jerusalem.

Liber II Paralipomenon

terea sacerdotum conbusit in altaribus idolorum mundauitque Iudam et Hierusalem 34:6 sed et in urbibus Manasse et 34:6 And so did he in the cities of Manasseh, and Ephraim,

and Simeon, even unto Naphtali, with their mattocks round about.

Ephraim et Symeon usque Nepthalim cuncta subuertit 34:7

34:7 And when he had broken down the altars and the groves, and had beaten the graven images into powder, and cut down all the idols throughout all the land of Israel, he returned to Jerusalem.

cumque altaria dissipasset et lucos et sculptilia contriuisset in frusta cunctaque delubra demolitus esset de uniuersa terra Israhel reuersus est Hierusalem 34:8 igitur anno octauo- 34:8 Now in the eighteenth year of his reign, when he had

purged the land, and the house, he sent Shaphan the son of Azaliah, and Maaseiah the governor of the city, and Joah the son of Joahaz the recorder, to repair the house of the LORD his God.

decimo regni sui mundata iam terra et templo Domini misit Saphan filium Eseliae et Maasiam principem ciuitatis et Ioha filium Ioachaz a commentariis ut instaurarent domum Domini Dei sui 34:9 qui uenerunt ad Helciam sacerdotem 34:9 And when they came to Hilkiah the high priest, they

delivered the money that was brought into the house of God, which the Levites that kept the doors had gathered of the hand of Manasseh and Ephraim, and of all the remnant of Israel, and of all Judah and Benjamin; and they returned to Jerusalem.

magnum acceptamque ab eo pecuniam quae inlata fuerat in domum Domini et quam congregauerant Leuitae ianitores de Manasse et Ephraim et uniuersis reliquiis Israhel ab omni quoque Iuda et Beniamin et habitatoribus Hierusalem 34:10 tradiderunt in manibus eorum qui praeerant operariis 34:10 And they put it in the hand of the workmen that had

the oversight of the house of the LORD, and they gave it to the workmen that wrought in the house of the LORD,

in domo Domini ut instaurarent templum et infirma quaeque to repair and amend the house: sarcirent 34:11 at illi dederunt eam artificibus et cementa- 34:11 Even to the artificers and builders gave they it, to

buy hewn stone, and timber for couplings, and to floor the houses which the kings of Judah had destroyed.

riis ut emerent lapides de lapidicinis et ligna ad commissuras aedificii et ad contignationem domorum quas destruxerant reges Iuda 34:12 qui fideliter cuncta faciebant erant autem 34:12 And the men did the work faithfully: and the over-

seers of them were Jahath and Obadiah, the Levites, of the sons of Merari; and Zechariah and Meshullam, of the sons of the Kohathites, to set it forward; and other of the Levites, all that could skill of instruments of musick.

praepositi operantium Iaath et Abdias de filiis Merari Zaccharias et Mosollam de filiis Caath qui urguebant opus omnes Leuitae scientes organis canere 34:13 super eos uero 34:13 Also they were over the bearers of burdens, and

were overseers of all that wrought the work in any manner of service: and of the Levites there were scribes, and

qui ad uarios usus onera portabant erant scribae et magistri officers, and porters. de Leuitis ianitores 34:14 cumque efferrent pecuniam quae 34:14 And when they brought out the money that was

brought into the house of the LORD, Hilkiah the priest found a book of the law of the LORD given by Moses.

inlata fuerat in templum Domini repperit Helcias sacerdos librum legis Domini per manum Mosi 34:15 et ait ad Sa- 34:15 And Hilkiah answered and said to Shaphan the

scribe, I have found the book of the law in the house of the LORD. And Hilkiah delivered the book to Shaphan.

phan scribam librum legis inueni in domo Domini et tradidit ei 34:16 at ille intulit uolumen ad regem et nuntiauit ei di- 34:16 And Shaphan carried the book to the king, and

brought the king word back again, saying, All that was committed to thy servants, they do it.

cens omnia quae dedisti in manu seruorum tuorum ecce con-

plentur 34:17 argentum quod reppertum est in domo Do- 34:17 And they have gathered together the money that was found in the house of the LORD, and have delivered it into the hand of the overseers, and to the hand of the workmen.

mini conflauerunt datumque est praefectis artificum et diuersa opera fabricantium 34:18 praeterea tradidit mihi Hel- 34:18 Then Shaphan the scribe told the king, saying,

Hilkiah the priest hath given me a book. And Shaphan read it before the king.

Biblia Sacra Vulgata

cias sacerdos hunc librum quem cum rege praesente recitas-

34:19 And it came to pass, when the king had heard the set 34:19 audissetque ille uerba legis scidit uestimenta sua

words of the law, that he rent his clothes.

34:20 And the king commanded Hilkiah, and Ahikam 34:20 et praecepit Helciae et Ahicam filio Saphan et Abdon

the son of Shaphan, and Abdon the son of Micah, and Shaphan the scribe, and Asaiah a servant of the king’s,

saying, filio Micha Saphan quoque scribae et Asaiae seruo regis di-34:21 Go, enquire of the LORD for me, and for them that cens 34:21 ite et orate Dominum pro me et pro reliquiis Is-

are left in Israel and in Judah, concerning the words of the book that is found: for great is the wrath of the LORD that is poured out upon us, because our fathers have not kept the word of the LORD, to do after all that is written in this book.

rahel et Iuda super uniuersis sermonibus libri istius qui reppertus est magnus enim furor Domini stillauit super nos eo quod non custodierint patres nostri uerba Domini ut face-

34:22 And Hilkiah, and they that the king had appointed, rent omnia quae scripta sunt in isto uolumine 34:22 abiit

went to Huldah the prophetess, the wife of Shallum the son of Tikvath, the son of Hasrah, keeper of the wardrobe; (now she dwelt in Jerusalem in the college:) and they spake to her to that effect.

igitur Helcias et hii qui simul a rege missi fuerant ad Holdan propheten uxorem Sellum filii Thecuath filii Hasra custodis uestium quae habitabat Hierusalem in secunda et locuti sunt

34:23 And she answered them, Thus saith the LORD God ei uerba quae supra narrauimus 34:23 at illa respondit eis

of Israel, Tell ye the man that sent you to me,

haec dicit Dominus Deus Israhel dicite uiro qui misit uos ad

34:24 Thus saith the LORD, Behold, I will bring evil upon me 34:24 haec dicit Dominus ecce ego inducam mala super this place, and upon the inhabitants thereof, even all the curses that are written in the book which they have read

before the king of Judah: locum istum et super habitatores eius cunctaque maledicta

quae scripta sunt in libro hoc quem legerunt coram rege Iuda

34:25 Because they have forsaken me, and have burned in- 34:25 quia dereliquerunt me et sacrificauerunt diis alienis ut

cense unto other gods, that they might provoke me to anger with all the works of their hands; therefore my wrath shall

be poured out upon this place, and shall not be quenched. me ad iracundiam prouocarent in cunctis operibus manuum

suarum idcirco stillauit furor meus super locum istum et non

34:26 And as for the king of Judah, who sent you to en- extinguetur 34:26 ad regem autem Iuda qui misit uos pro

quire of the LORD, so shall ye say unto him, Thus saith the LORD God of Israel concerning the words which thou

hast heard; Domino deprecando sic loquimini haec dicit Dominus Deus 34:27 Because thine heart was tender, and thou didst hum- Israhel quoniam audisti uerba uoluminis 34:27 atque emol-

ble thyself before God, when thou heardest his words against this place, and against the inhabitants thereof, and humbledst thyself before me, and didst rend thy clothes, and weep before me; I have even heard thee also, saith the LORD.

litum est cor tuum et humiliatus es in conspectu Dei super his quae dicta sunt contra locum hunc et habitatores Hierusalem reueritusque faciem meam scidisti uestimenta tua et fleuisti coram me ego quoque exaudiui te dicit Domi-

34:28 Behold, I will gather thee to thy fathers, and thou nus 34:28 iam enim colligam te ad patres tuos et infereshalt be gathered to thy grave in peace, neither shall thine eyes see all the evil that I will bring upon this place, and upon the inhabitants of the same. So they brought the king word again.

ris in sepulchrum tuum in pace nec uidebunt oculi tui omne malum quod ego inducturus sum super locum istum et super habitatores eius rettulerunt itaque regi cuncta quae dixe-

34:29 Then the king sent and gathered together all the el- rat 34:29 at ille conuocatis uniuersis maioribus natu Iuda et

ders of Judah and Jerusalem.

34:30 And the king went up into the house of the LORD, Hierusalem 34:30 ascendit domum Domini unaque omnes

and all the men of Judah, and the inhabitants of Jerusalem, and the priests, and the Levites, and all the people, great and small: and he read in their ears all the words of the book of the covenant that was found in the house of the LORD.

Liber II Paralipomenon

uiri Iuda et habitatores Hierusalem sacerdotes et Leuitae et cunctus populus a minimo usque ad maximum quibus audientibus in domo Domini legit rex omnia uerba uoluminis 34:31 et stans in tribunali suo percussit foedus coram Domino 34:31 And the king stood in his place, and made a

covenant before the LORD, to walk after the LORD, and to keep his commandments, and his testimonies, and his statutes, with all his heart, and with all his soul, to perform the words of the covenant which are written in this book.

ut ambularet post eum et custodiret praecepta et testimonia et iustificationes eius in toto corde suo et in tota anima sua faceretque quae scripta sunt in uolumine illo quem legerat 34:32 adiurauit quoque super hoc omnes qui repperti fuerant 34:32 And he caused all that were present in Jerusalem and

Benjamin to stand to it. And the inhabitants of Jerusalem did according to the covenant of God, the God of their

in Hierusalem et Beniamin et fecerunt habitatores Hierusa- fathers. lem iuxta pactum Domini Dei patrum suorum 34:33 abstu- 34:33 And Josiah took away all the abominations out of

all the countries that pertained to the children of Israel, and made all that were present in Israel to serve, even to serve the LORD their God. And all his days they departed not from following the LORD, the God of their fathers.

lit ergo Iosias cunctas abominationes de uniuersis regionibus filiorum Israhel et fecit omnes qui residui erant in Israhel seruire Domino Deo suo cunctis diebus eius non recesserunt a Domino Deo patrum suorum

35:1 fecit autem Iosias in Hierusalem phase Domino quod 35:1 Moreover Josiah kept a passover unto the LORD in

Jerusalem: and they killed the passover on the fourteenth day of the first month.

immolatum est quartadecima die mensis primi 35:2 et con-

35:2 And he set the priests in their charges, and encouraged them to the service of the house of the LORD,

stituit sacerdotes in officiis suis hortatusque est eos ut mi-

nistrarent in domo Domini 35:3 Leuitis quoque ad quorum 35:3 And said unto the Levites that taught all Israel, which were holy unto the LORD, Put the holy ark in the house which Solomon the son of David king of Israel did build; it shall not be a burden upon your shoulders: serve now the LORD your God, and his people Israel,

eruditionem omnis Israhel sanctificabatur Domino locutus est ponite arcam in sanctuario templi quod aedificauit Salomon filius Dauid rex Israhel nequaquam enim eam ultra portabitis nunc autem ministrate Domino Deo uestro et populo eius Israhel 35:4 et praeparate uos per domos et cognatio- 35:4 And prepare yourselves by the houses of your fathers,

after your courses, according to the writing of David king of Israel, and according to the writing of Solomon his son.

nes uestras in diuisionibus singulorum sicut praecepit Dauid rex Israhel et descripsit Salomon filius eius 35:5 ministrate 35:5 And stand in the holy place according to the divisions

of the families of the fathers of your brethren the people, and after the division of the families of the Levites.

in sanctuario per familias turmasque leuiticas 35:6 et sanc-

35:6 So kill the passover, and sanctify yourselves, and prepare your brethren, that they may do according to the word of the LORD by the hand of Moses.

tificati immolate phase fratres etiam uestros ut possint iuxta uerba quae locutus est Dominus in manu Mosi facere praeparate 35:7 dedit praeterea Iosias omni populo qui ibi fue- 35:7 And Josiah gave to the people, of the flock, lambs

and kids, all for the passover offerings, for all that were present, to the number of thirty thousand, and three thou-

rat inuentus in sollemnitatem phase agnos et hedos de gregi- sand bullocks: these were of the king’s substance. bus et reliqui pecoris triginta milia boumque tria milia haec de regis uniuersa substantia 35:8 duces quoque eius sponte 35:8 And his princes gave willingly unto the people, to

the priests, and to the Levites: Hilkiah and Zechariah and Jehiel, rulers of the house of God, gave unto the priests for the passover offerings two thousand and six hundred small cattle and three hundred oxen.

quod uoluerant obtulerunt tam populo quam sacerdotibus et

Biblia Sacra Vulgata

Leuitis porro Helcias et Zaccharias et Iehihel principes domus Domini dederunt sacerdotibus ad faciendum phase pe-

35:9 Conaniah also, and Shemaiah and Nethaneel, his cora commixtim duo milia sescenta et boues trecentos 35:9

brethren, and Hashabiah and Jeiel and Jozabad, chief of the Levites, gave unto the Levites for passover offerings

five thousand small cattle, and five hundred oxen. Chonenias autem Semeias etiam et Nathanahel fratres eius necnon Asabias et Iahihel et Iozabath principes Leuitarum dederunt ceteris Leuitis ad celebrandum phase quinque milia

35:10 So the service was prepared, and the priests stood in pecorum et boues quingentos 35:10 praeparatumque est mitheir place, and the Levites in their courses, according to the king’s commandment.

nisterium et steterunt sacerdotes in officio suo Leuitae quo-

35:11 And they killed the passover, and the priests sprin- que in turmis iuxta regis imperium 35:11 et immolatum est kled the blood from their hands, and the Levites flayed them.

phase asperseruntque sacerdotes manu sua sanguinem et Le-

35:12 And they removed the burnt offerings, that they uitae detraxerunt pelles holocaustorum 35:12 et separaue-

might give according to the divisions of the families of the people, to offer unto the LORD, as it is written in the

book of Moses. And so did they with the oxen. runt ea ut darent per domos et familias singulorum et of-

ferrentur Domino sicut scriptum est in libro Mosi de bubus

35:13 And they roasted the passover with fire according to quoque fecere similiter 35:13 et assauerunt phase super ig-

the ordinance: but the other holy offerings sod they in pots, and in caldrons, and in pans, and divided them speedily

among all the people. nem iuxta quod lege praeceptum est pacificas uero hostias

coxerunt in lebetis et caccabis et ollis et festinato distribue-

35:14 And afterward they made ready for themselves, and runt uniuersae plebi 35:14 sibi autem et sacerdotibus postea

for the priests: because the priests the sons of Aaron were busied in offering of burnt offerings and the fat until night; therefore the Levites prepared for themselves, and for the priests the sons of Aaron.

parauerunt nam in oblatione holocaustorum et adipum usque ad noctem sacerdotes fuerant occupati unde Leuitae et sibi et

35:15 And the singers the sons of Asaph were in their sacerdotibus filiis Aaron parauerunt nouissimis 35:15 porro

place, according to the commandment of David, and Asaph, and Heman, and Jeduthun the king’s seer; and the porters waited at every gate; they might not depart from their service; for their brethren the Levites prepared for them.

cantores filii Asaph stabant in ordine suo iuxta praeceptum Dauid et Asaph et Heman et Idithun prophetarum regis ianitores uero per portas singulas obseruabant ita ut ne puncto quidem discederent a ministerio quam ob rem et fratres eo-

35:16 So all the service of the LORD was prepared the rum Leuitae parauerunt eis cibos 35:16 omnis igitur cul-

same day, to keep the passover, and to offer burnt offerings upon the altar of the LORD, according to the command-

ment of king Josiah. tura Domini rite conpleta est in die illa ut facerent phase et

offerrent holocausta super altare Domini iuxta praeceptum

35:17 And the children of Israel that were present kept the regis Iosiae 35:17 feceruntque filii Israhel qui repperti fue-

passover at that time, and the feast of unleavened bread seven days.

rant ibi phase in tempore illo et sollemnitatem azymorum

35:18 And there was no passover like to that kept in Israel septem diebus 35:18 non fuit phase simile huic in Israhel from the days of Samuel the prophet; neither did all the kings of Israel keep such a passover as Josiah kept, and the priests, and the Levites, and all Judah and Israel that were present, and the inhabitants of Jerusalem.

a diebus Samuhelis prophetae sed nec quisquam de cunctis regibus Israhel fecit phase sicut Iosias sacerdotibus et Leuitis et omni Iuda et Israhel qui reppertus fuerat et habitan-

Liber II Paralipomenon

tibus in Hierusalem 35:19 octauodecimo anno regni Iosiae 35:19 In the eighteenth year of the reign of Josiah was this

passover kept.

hoc phase celebratum est 35:20 postquam instaurauerat Io- 35:20 After all this, when Josiah had prepared the temple,

Necho king of Egypt came up to fight against Charchemish by Euphrates: and Josiah went out against him.

sias templum ascendit Nechao rex Aegypti ad pugnandum in Charchamis iuxta Eufraten et processit in occursum eius Iosias 35:21 at ille missis ad eum nuntiis ait quid mihi et tibi 35:21 But he sent ambassadors to him, saying, What have

I to do with thee, thou king of Judah? I come not against thee this day, but against the house wherewith I have war: for God commanded me to make haste: forbear thee from meddling with God, who is with me, that he destroy thee not.

est rex Iuda non aduersum te hodie uenio sed contra aliam pugno domum ad quam me Deus festinato ire praecepit desine aduersum Deum facere qui mecum est ne interficiat te 35:22 noluit Iosias reuerti sed praeparauit contra eum bellum 35:22 Nevertheless Josiah would not turn his face from

him, but disguised himself, that he might fight with him, and hearkened not unto the words of Necho from the

nec adquieuit sermonibus Nechao ex ore Dei uerum perrexit mouth of God, and came to fight in the valley of Megiddo. ut dimicaret in campo Mageddo 35:23 ibique uulneratus a 35:23 And the archers shot at king Josiah; and the king

said to his servants, Have me away; for I am sore wounded.

sagittariis dixit pueris suis educite me de proelio quia op-

pido uulneratus sum 35:24 qui transtulerunt eum de curru 35:24 His servants therefore took him out of that chariot, and put him in the second chariot that he had; and they brought him to Jerusalem, and he died, and was buried in one of the sepulchres of his fathers. And all Judah and Jerusalem mourned for Josiah.

in alterum currum qui sequebatur eum more regio et asportauerunt in Hierusalem mortuusque est et sepultus in mausoleo patrum suorum et uniuersus Iuda et Hierusalem luxerunt eum 35:25 Hieremias maxime cuius omnes cantores atque 35:25 And Jeremiah lamented for Josiah: and all the

singing men and the singing women spake of Josiah in their lamentations to this day, and made them an ordinance

cantrices usque in praesentem diem lamentationes super Io- in Israel: and, behold, they are written in the lamentations.

sia replicant et quasi lex obtinuit in Israhel ecce scriptum fertur in Lamentationibus 35:26 reliqua autem sermonum 35:26 Now the rest of the acts of Josiah, and his good-

ness, according to that which was written in the law of the LORD,

Iosiae et misericordiarum eius quae lege praecepta sunt Do-

mini 35:27 opera quoque illius prima et nouissima scripta 35:27 And his deeds, first and last, behold, they are written

in the book of the kings of Israel and Judah.

sunt in libro regum Israhel et Iuda

36:1 tulit ergo populus terrae Ioachaz filium Iosiae et con- 36:1 Then the people of the land took Jehoahaz the son

of Josiah, and made him king in his father’s stead in Jerusalem.

stituit regem pro patre suo in Hierusalem 36:2 uiginti trium

36:2 Jehoahaz was twenty and three years old when he began to reign, and he reigned three months in Jerusalem.

annorum erat Ioachaz cum regnare coepisset et tribus mensibus regnauit in Hierusalem 36:3 amouit autem eum rex Ae- 36:3 And the king of Egypt put him down at Jerusalem,

and condemned the land in an hundred talents of silver and a talent of gold.

gypti cum uenisset Hierusalem et condemnauit terram centum talentis argenti et talento auri 36:4 constituitque regem 36:4 And the king of Egypt made Eliakim his brother

king over Judah and Jerusalem, and turned his name to Jehoiakim. And Necho took Jehoahaz his brother, and car-

pro eo Eliacim fratrem eius super Iudam et Hierusalem et ried him to Egypt. uertit nomen eius Ioacim ipsum uero Ioachaz tulit secum et adduxit in Aegyptum 36:5 uiginti quinque annorum erat 36:5 Jehoiakim was twenty and five years old when he

began to reign, and he reigned eleven years in Jerusalem: and he did that which was evil in the sight of the LORD

Ioacim cum regnare coepisset et undecim annis regnauit in his God.

Biblia Sacra Vulgata

36:6 Against him came up Nebuchadnezzar king of Baby- Hierusalem fecitque malum coram Domino Deo suo 36:6 lon, and bound him in fetters, to carry him to Babylon.

contra hunc ascendit Nabuchodonosor rex Chaldeorum et

36:7 Nebuchadnezzar also carried of the vessels of the uinctum catenis duxit in Babylonem 36:7 ad quam et uasa house of the LORD to Babylon, and put them in his temple at Babylon.

Domini transtulit et posuit ea in templo suo 36:8 reliqua

36:8 Now the rest of the acts of Jehoiakim, and his abominations which he did, and that which was found in him, behold, they are written in the book of the kings of Israel and Judah: and Jehoiachin his son reigned in his stead.

autem uerborum Ioacim et abominationum eius quas operatus est et quae inuenta sunt in eo continentur in libro regum Israhel et Iuda regnauitque Ioachin filius eius pro eo

36:9 Jehoiachin was eight years old when he began to 36:9 octo annorum erat Ioachin cum regnare coepisset et tri-

reign, and he reigned three months and ten days in Jerusalem: and he did that which was evil in the sight of

the LORD. bus mensibus ac decem diebus regnauit in Hierusalem fe-36:10 And when the year was expired, king Nebuchad- citque malum in conspectu Domini 36:10 cumque anni cir-

nezzar sent, and brought him to Babylon, with the goodly vessels of the house of the LORD, and made Zedekiah his

brother king over Judah and Jerusalem. culus uolueretur misit Nabuchodonosor rex qui et adduxerunt eum in Babylonem asportatis simul pretiosissimis uasis domus Domini regem uero constituit Sedeciam fratrem

36:11 Zedekiah was one and twenty years old when he eius super Iudam et Hierusalem 36:11 uiginti et unius anni

began to reign, and reigned eleven years in Jerusalem.

erat Sedecias cum regnare coepisset et undecim annis regna-

36:12 And he did that which was evil in the sight of the uit in Hierusalem 36:12 fecitque malum in oculis Domini

LORD his God, and humbled not himself before Jeremiah the prophet speaking from the mouth of the LORD.

Dei sui nec erubuit faciem Hieremiae prophetae loquentis

36:13 And he also rebelled against king Nebuchadnezzar, ad se ex ore Domini 36:13 a rege quoque Nabuchodonowho had made him swear by God: but he stiffened his neck, and hardened his heart from turning unto the LORD

God of Israel. sor recessit qui adiurauerat eum per Deum et indurauit cer-

uicem suam et cor ut non reuerteretur ad Dominum Deum

36:14 Moreover all the chief of the priests, and the people, Israhel 36:14 sed et uniuersi principes sacerdotum et popu-

transgressed very much after all the abominations of the heathen; and polluted the house of the LORD which he

had hallowed in Jerusalem. lus praeuaricati sunt inique iuxta uniuersas abominationes

gentium et polluerunt domum Domini quam sanctificauerat

36:15 And the LORD God of their fathers sent to them by sibi in Hierusalem 36:15 mittebat autem Dominus Deus pa-

his messengers, rising up betimes, and sending; because he had compassion on his people, and on his dwelling

place: trum suorum ad illos per manum nuntiorum suorum de nocte

consurgens et cotidie commonens eo quod parceret populo

36:16 But they mocked the messengers of God, and de- et habitaculo suo 36:16 at illi subsannabant nuntios Dei et

spised his words, and misused his prophets, until the wrath of the LORD arose against his people, till there was no

remedy. paruipendebant sermones eius inludebantque prophetis do-

nec ascenderet furor Domini in populum eius et esset nulla

36:17 Therefore he brought upon them the king of the curatio 36:17 adduxit enim super eos regem Chaldeorum et

Chaldees, who slew their young men with the sword in the house of their sanctuary, and had no compassion upon young man or maiden, old man, or him that stooped for age: he gave them all into his hand.

interfecit iuuenes eorum gladio in domo sanctuarii sui non est misertus adulescentis et uirginis et senis nec decrepiti

36:18 And all the vessels of the house of God, great and quidem sed omnes tradidit manibus eius 36:18 uniuersaque

small, and the treasures of the house of the LORD, and the treasures of the king, and of his princes; all these he brought to Babylon.

Liber II Paralipomenon

uasa domus Domini tam maiora quam minora et thesauros templi et regis et principum transtulit in Babylonem 36:19 36:19 And they burnt the house of God, and brake down

the wall of Jerusalem, and burnt all the palaces thereof with fire, and destroyed all the goodly vessels thereof.

incenderunt hostes domum Dei destruxerunt murum Hierusalem uniuersas turres conbuserunt et quicquid pretiosum fuerat demoliti sunt 36:20 si quis euaserat gladium ductus 36:20 And them that had escaped from the sword carried

he away to Babylon; where they were servants to him and his sons until the reign of the kingdom of Persia:

in Babylonem seruiuit regi et filiis eius donec imperaret rex Persarum 36:21 et conpleretur sermo Domini ex ore Hie- 36:21 To fulfil the word of the LORD by the mouth of

Jeremiah, until the land had enjoyed her sabbaths: for as long as she lay desolate she kept sabbath, to fulfil three-

remiae et celebraret terra sabbata sua cunctis enim diebus score and ten years. desolationis egit sabbatum usque dum conplerentur septua-

ginta anni 36:22 anno autem primo Cyri regis Persarum ad 36:22 Now in the first year of Cyrus king of Persia, that the word of the LORD spoken by the mouth of Jeremiah might be accomplished, the LORD stirred up the spirit of Cyrus king of Persia, that he made a proclamation throughout all his kingdom, and put it also in writing, saying,

explendum sermonem Domini quem locutus fuerat per os Hieremiae suscitauit Dominus spiritum Cyri regis Persarum qui iussit praedicari in uniuerso regno suo etiam per scri-

pturam dicens 36:23 haec dicit Cyrus rex Persarum omnia 36:23 Thus saith Cyrus king of Persia, All the kingdoms of the earth hath the LORD God of heaven given me; and he hath charged me to build him an house in Jerusalem, which is in Judah. Who is there among you of all his people? The LORD his God be with him, and let him go up.

regna terrae dedit mihi Dominus Deus caeli et ipse praecepit mihi ut aedificarem ei domum in Hierusalem quae est in Iudaea quis ex uobis est in omni populo eius sit Dominus Deus suus cum eo et ascendat

Biblia Sacra Vulgata

DCCXCVI

\underline{Liber Esdrae}

1:1 in anno primo Cyri regis Persarum ut conpleretur uer- 1:1 Now in the first year of Cyrus king of Persia, that

the word of the LORD by the mouth of Jeremiah might be fulfilled, the LORD stirred up the spirit of Cyrus king of Persia, that he made a proclamation throughout all his kingdom, and put it also in writing, saying,

bum Domini ex ore Hieremiae suscitauit Dominus spiritum Cyri regis Persarum et transduxit uocem in uniuerso regno suo etiam per scripturam dicens 1:2 haec dicit Cyrus rex 1:2 Thus saith Cyrus king of Persia, The LORD God of

heaven hath given me all the kingdoms of the earth; and he hath charged me to build him an house at Jerusalem,

Persarum omnia regna terrae dedit mihi Dominus Deus caeli which is in Judah. et ipse praecepit mihi ut aedificarem ei domum in Hierusalem quae est in Iudaea 1:3 quis est in uobis de uniuerso po- 1:3 Who is there among you of all his people? his God

be with him, and let him go up to Jerusalem, which is in Judah, and build the house of the LORD God of Israel, (he

pulo eius sit Deus illius cum ipso ascendat Hierusalem quae is the God,) which is in Jerusalem. est in Iudaea et aedificet domum Domini Dei Israhel ipse est Deus qui est in Hierusalem 1:4 et omnes reliqui in cunctis 1:4 And whosoever remaineth in any place where he so-

journeth, let the men of his place help him with silver, and with gold, and with goods, and with beasts, beside the

locis ubicumque habitant adiuuent eum uiri de loco suo ar- freewill offering for the house of God that is in Jerusalem. gento et auro et substantia et pecoribus excepto quod uoluntarie offerunt templo Dei quod est in Hierusalem 1:5 et sur- 1:5 Then rose up the chief of the fathers of Judah and

Benjamin, and the priests, and the Levites, with all them whose spirit God had raised, to go up to build the house of

rexerunt principes patrum de Iuda et Beniamin et sacerdotes the LORD which is in Jerusalem. et Leuitae omnis cuius suscitauit Deus spiritum ut ascenderent ad aedificandum templum Domini quod erat in Hierusalem 1:6 uniuersique qui erant in circuitu adiuuerunt manus 1:6 And all they that were about them strengthened their

hands with vessels of silver, with gold, with goods, and with beasts, and with precious things, beside all that was

eorum in uasis argenteis et aureis in substantia in iumentis willingly offered. in supellectili exceptis his quae sponte obtulerunt 1:7 rex 1:7 Also Cyrus the king brought forth the vessels of the

house of the LORD, which Nebuchadnezzar had brought forth out of Jerusalem, and had put them in the house of

quoque Cyrus protulit uasa templi Domini quae tulerat Na- his gods; buchodonosor de Hierusalem et posuerat ea in templo dei sui 1:8 protulit autem ea Cyrus rex Persarum per manum Mitri- 1:8 Even those did Cyrus king of Persia bring forth by the

hand of Mithredath the treasurer, and numbered them unto Sheshbazzar, the prince of Judah.

datis filii Gazabar et adnumerauit ea Sasabassar principi Iu-

dae 1:9 et hic est numerus eorum fialae aureae triginta fia- 1:9 And this is the number of them: thirty chargers of gold,

a thousand chargers of silver, nine and twenty knives,

lae argenteae mille cultri uiginti nouem scyphi aurei triginta 1:10 scyphi argentei secundi quadringenti decem uasa alia 1:10 Thirty basons of gold, silver basons of a second sort

four hundred and ten, and other vessels a thousand.

mille 1:11 omnia uasa aurea et argentea quinque milia qua- 1:11 All the vessels of gold and of silver were five thou-

sand and four hundred. All these did Sheshbazzar bring up with them of the captivity that were brought up from

dringenta uniuersa tulit Sasabassar cum his qui ascendebant Babylon unto Jerusalem. de transmigratione Babylonis in Hierusalem

2:1 hii sunt autem filii prouinciae qui ascenderunt de ca- 2:1 Now these are the children of the province that went up

out of the captivity, of those which had been carried away, whom Nebuchadnezzar the king of Babylon had carried away unto Babylon, and came again unto Jerusalem and Judah, every one unto his city;

ptiuitate quam transtulerat Nabuchodonosor rex Babylonis in Babylonem et reuersi sunt in Hierusalem et Iudam unusquisque in ciuitatem suam 2:2 qui uenerunt cum Zorobabel 2:2 Which came with Zerubbabel: Jeshua, Nehemiah,

Seraiah, Reelaiah, Mordecai, Bilshan, Mizpar, Bigvai, Rehum, Baanah. The number of the men of the people of Israel:

Hiesua Neemia Saraia Rahelaia Mardochai Belsan Mesphar

2:3 The children of Parosh, two thousand an hundred sev- Beguai Reum Baana numerus uirorum populi Israhel 2:3 enty and two. 2:4 The children of Shephatiah, three hundred seventy and filii Pharos duo milia centum septuaginta duo 2:4 filii Setwo.

2:5 The children of Arah, seven hundred seventy and five. phetia trecenti septuaginta duo 2:5 filii Area septingenti se-2:6 The children of Pahathmoab, of the children of Jeshua ptuaginta quinque 2:6 filii Phaethmoab filiorum Iosue Ioab

and Joab, two thousand eight hundred and twelve.

2:7 The children of Elam, a thousand two hundred fifty duo milia octingenti duodecim 2:7 filii Helam mille ducenti

and four.

2:8 The children of Zattu, nine hundred forty and five. quinquaginta quattuor 2:8 filii Zeththua nongenti quadra-2:9 The children of Zaccai, seven hundred and threescore. ginta quinque 2:9 filii Zacchai septingenti sexaginta 2:10

2:10 The children of Bani, six hundred forty and two.

2:11 The children of Bebai, six hundred twenty and three. filii Bani sescenti quadraginta duo 2:11 filii Bebai sescenti 2:12 The children of Azgad, a thousand two hundred uiginti tres 2:12 filii Azgad mille ducenti uiginti duo 2:13

twenty and two. 2:13 The children of Adonikam, six hundred sixty and six.

filii Adonicam sescenti sexaginta sex 2:14 filii Beguai duo

2:14 The children of Bigvai, two thousand fifty and six.

2:15 The children of Adin, four hundred fifty and four. milia quinquaginta sex 2:15 filii Adin quadringenti quin-2:16 The children of Ater of Hezekiah, ninety and eight. quaginta quattuor 2:16 filii Ater qui erant ex Hiezechia no-2:17 The children of Bezai, three hundred twenty and naginta octo 2:17 filii Besai trecenti uiginti tres 2:18 filii

three. 2:18 The children of Jorah, an hundred and twelve.

Iora centum duodecim 2:19 filii Asom ducenti uiginti tres

2:19 The children of Hashum, two hundred twenty and three.

2:20 filii Gebbar nonaginta quinque 2:21 filii Bethleem cen-

2:20 The children of Gibbar, ninety and five. 2:21 The children of Bethlehem, an hundred twenty and

three. tum uiginti tres 2:22 uiri Netupha quinquaginta sex 2:23

2:22 The men of Netophah, fifty and six.

2:23 The men of Anathoth, an hundred twenty and eight. uiri Anathoth centum uiginti octo 2:24 filii Azmaueth qua-

2:24 The children of Azmaveth, forty and two. 2:25 The children of Kirjatharim, Chephirah, and Beeroth, draginta duo 2:25 filii Cariathiarim Caephira et Beroth seseven hundred and forty and three.

2:26 The children of Ramah and Gaba, six hundred twenty ptingenti quadraginta tres 2:26 filii Arama et Gaba sescenti

and one.

2:27 The men of Michmas, an hundred twenty and two. uiginti unus 2:27 uiri Machmas centum uiginti duo 2:28

2:28 The men of Bethel and Ai, two hundred twenty and three.

uiri Bethel et Gai ducenti uiginti tres 2:29 filii Nebo quin-

2:29 The children of Nebo, fifty and two.

2:30 The children of Magbish, an hundred fifty and six. quaginta duo 2:30 filii Megbis centum quinquaginta sex 2:31 The children of the other Elam, a thousand two hun- 2:31 filii Helam alterius mille ducenti quinquaginta quattuor

dred fifty and four.

2:32 The children of Harim, three hundred and twenty. 2:32 filii Arim trecenti uiginti 2:33 filii Lod Adid et Ono se-

2:33 The children of Lod, Hadid, and Ono, seven hundred twenty and five.

ptingenti uiginti quinque 2:34 filii Hiericho trecenti quad-

2:34 The children of Jericho, three hundred forty and five.

2:35 The children of Senaah, three thousand and six hun- raginta quinque 2:35 filii Sennaa tria milia sescenti triginta

dred and thirty.

2:36 The priests: the children of Jedaiah, of the house of 2:36 sacerdotes filii Idaia in domo Hiesue nongenti septua-

Jeshua, nine hundred seventy and three.

2:37 The children of Immer, a thousand fifty and two. ginta tres 2:37 filii Emmer mille quinquaginta duo 2:38 fi-

2:38 The children of Pashur, a thousand two hundred forty and seven.

lii Phessur mille ducenti quadraginta septem 2:39 filii Arim

2:39 The children of Harim, a thousand and seventeen. 2:40 The Levites: the children of Jeshua and Kadmiel, of mille decem et septem 2:40 Leuitae filii Hiesue et Cedmithe children of Hodaviah, seventy and four.

2:41 The singers: the children of Asaph, an hundred hel filiorum Odeuia septuaginta quattuor 2:41 cantores filii

twenty and eight. 2:42 The children of the porters: the children of Shallum, Asaph centum uiginti octo 2:42 filii ianitorum filii Sellum the children of Ater, the children of Talmon, the children of Akkub, the children of Hatita, the children of Shobai, in all an hundred thirty and nine.

filii Ater filii Telmon filii Accub filii Atita filii Sobai uniuersi centum triginta nouem 2:43 Nathinnei filii Sia filii Asupha 2:43 The Nethinims: the children of Ziha, the children of

Hasupha, the children of Tabbaoth,

filii Tebbaoth 2:44 filii Ceros filii Siaa filii Phadon 2:45 2:44 The children of Keros, the children of Siaha, the children of Padon,

2:45 The children of Lebanah, the children of Hagabah, the children of Akkub,

filii Leuana filii Agaba filii Accub 2:46 filii Agab filii Sel-

2:46 The children of Hagab, the children of Shalmai, the

mai filii Anan 2:47 filii Gaddel filii Gaer filii Rahaia 2:48 _{children of Hanan,}

2:47 The children of Giddel, the children of Gahar, the children of Reaiah,

2:48 The children of Rezin, the children of Nekoda, the children of Gazzam,

filii Rasin filii Nechoda filii Gazem 2:49 filii Aza filii Pha-

2:49 The children of Uzza, the children of Paseah, the children of Besai,

sea filii Besee 2:50 filii Asenaa filii Munim filii Nephusim

2:50 The children of Asnah, the children of Mehunim, the children of Nephusim,

2:51 filii Becbuc filii Acupha filii Arur 2:52 filii Besluth fi-

2:51 The children of Bakbuk, the children of Hakupha, the children of Harhur,

2:52 The children of Bazluth, the children of Mehida, the children of Harsha,

lii Maida filii Arsa 2:53 filii Bercos filii Sisara filii Thema

2:53 The children of Barkos, the children of Sisera, the children of Thamah,

2:54 filii Nasia filii Atupha 2:55 filii seruorum Salomonis

2:54 The children of Neziah, the children of Hatipha.

2:55 The children of Solomon’s servants: the children of Sotai, the children of Sophereth, the children of Peruda,

filii Sotei filii Suphereth filii Pharuda 2:56 filii Iala filii De-

2:56 The children of Jaalah, the children of Darkon, the children of Giddel,

rcon filii Gedel 2:57 filii Saphatia filii Athil filii Phocereth

2:57 The children of Shephatiah, the children of Hattil, the children of Pochereth of Zebaim, the children of Ami.

qui erant de Asebaim filii Ammi 2:58 omnes Nathinnei et

2:58 All the Nethinims, and the children of Solomon’s servants, were three hundred ninety and two.

filii seruorum Salomonis trecenti nonaginta duo 2:59 et hii

2:59 And these were they which went up from Telmelah, Telharsa, Cherub, Addan, and Immer: but they could not shew their father’s house, and their seed, whether they were of Israel:

qui ascenderunt de Thelmela Thelarsa Cherub et Don et Mer et non potuerunt indicare domum patrum suorum et semen suum utrum ex Israhel essent 2:60 filii Delaia filii Tobia filii 2:60 The children of Delaiah, the children of Tobiah, the

children of Nekoda, six hundred fifty and two.

Necoda sescenti quinquaginta duo 2:61 et de filiis sacerdo- 2:61 And of the children of the priests: the children of

Habaiah, the children of Koz, the children of Barzillai; which took a wife of the daughters of Barzillai the Gilea-

tum filii Obia filii Accos filii Berzellai qui accepit de filia- dite, and was called after their name: bus Berzellai Galaditis uxorem et uocatus est nomine eorum 2:62 hii quaesierunt scripturam genealogiae suae et non inue- 2:62 These sought their register among those that were

reckoned by genealogy, but they were not found: therefore were they, as polluted, put from the priesthood.

nerunt et eiecti sunt de sacerdotio 2:63 et dixit Athersatha

2:63 And the Tirshatha said unto them, that they should not eat of the most holy things, till there stood up a priest with Urim and with Thummim.

eis ut non comederent de sancto sanctorum donec surgeret sacerdos doctus atque perfectus 2:64 omnis multitudo quasi 2:64 The whole congregation together was forty and two

thousand three hundred and threescore,

unus quadraginta duo milia trecenti sexaginta 2:65 excep- 2:65 Beside their servants and their maids, of whom there

were seven thousand three hundred thirty and seven: and there were among them two hundred singing men and

tis seruis eorum et ancillis qui erant septem milia trecenti singing women. triginta septem et in ipsis cantores atque cantrices ducentae 2:66 equi eorum septingenti triginta sex muli eorum ducenti 2:66 Their horses were seven hundred thirty and six; their

mules, two hundred forty and five;

quadraginta quinque 2:67 cameli eorum quadringenti tri- 2:67 Their camels, four hundred thirty and five; their

asses, six thousand seven hundred and twenty.

ginta quinque asini eorum sex milia septingenti uiginti 2:68 2:68 And some of the chief of the fathers, when they came

to the house of the LORD which is at Jerusalem, offered freely for the house of God to set it up in his place:

et de principibus patrum cum ingrederentur templum Domini quod est in Hierusalem sponte obtulerunt in domum

Dei ad extruendam eam in loco suo 2:69 secundum uires 2:69 They gave after their ability unto the treasure of the work threescore and one thousand drams of gold, and five thousand pound of silver, and one hundred priests’ garments.

Biblia Sacra Vulgata

suas dederunt in inpensas operis auri solidos sexaginta milia et mille argenti minas quinque milia et uestes sacerdotales

2:70 So the priests, and the Levites, and some of the peo- centum 2:70 habitauerunt ergo sacerdotes et Leuitae et de

ple, and the singers, and the porters, and the Nethinims, dwelt in their cities, and all Israel in their cities.

populo et cantores et ianitores et Nathinnei in urbibus suis uniuersusque Israhel in ciuitatibus suis

3:1 And when the seventh month was come, and the chil- 3:1 iamque uenerat mensis septimus et erant filii Israhel in

dren of Israel were in the cities, the people gathered themselves together as one man to Jerusalem.

ciuitatibus suis congregatus est ergo populus quasi uir unus

3:2 Then stood up Jeshua the son of Jozadak, and his in Hierusalem 3:2 et surrexit Iosue filius Iosedech et fratres

brethren the priests, and Zerubbabel the son of Shealtiel, and his brethren, and builded the altar of the God of Israel, to offer burnt offerings thereon, as it is written in the law of Moses the man of God.

eius sacerdotes et Zorobabel filius Salathihel et fratres eius et aedificauerunt altare Dei Israhel ut offerrent in eo holocau-

3:3 And they set the altar upon his bases; for fear was stomata sicut scriptum est in lege Mosi uiri Dei 3:3 conlo-

upon them because of the people of those countries: and they offered burnt offerings thereon unto the LORD, even

burnt offerings morning and evening. cauerunt autem altare super bases suas deterrentibus eos per

circuitum populis terrarum et obtulerunt super illud holocau-

3:4 They kept also the feast of tabernacles, as it is written, stum Domino mane et uespere 3:4 feceruntque sollemnitaand offered the daily burnt offerings by number, according to the custom, as the duty of every day required;

tem tabernaculorum sicut scriptum est et holocaustum diebus singulis per ordinem secundum praeceptum opus diei in

3:5 And afterward offered the continual burnt offering, die suo 3:5 et post haec holocaustum iuge tam in kalendis

both of the new moons, and of all the set feasts of the LORD that were consecrated, and of every one that will-

ingly offered a freewill offering unto the LORD. quam in uniuersis sollemnitatibus Domini quae erant conse-

cratae et in omnibus in quibus ultro offerebatur munus Deo

3:6 From the first day of the seventh month began they to 3:6 a primo die mensis septimi coeperunt offerre holocaus-

offer burnt offerings unto the LORD. But the foundation of the temple of the LORD was not yet laid.

tum Domino porro templum Dei fundatum necdum erat 3:7

3:7 They gave money also unto the masons, and to the carpenters; and meat, and drink, and oil, unto them of Zidon, and to them of Tyre, to bring cedar trees from Lebanon to the sea of Joppa, according to the grant that they had of Cyrus king of Persia.

dederunt autem pecunias latomis et cementariis cibum quoque et potum et oleum Sidoniis Tyriisque ut deferrent ligna cedrina de Libano ad mare Ioppes iuxta quod praeceperat

3:8 Now in the second year of their coming unto the house Cyrus rex Persarum eis 3:8 anno autem secundo aduentus of God at Jerusalem, in the second month, began Zerubbabel the son of Shealtiel, and Jeshua the son of Jozadak, and the remnant of their brethren the priests and the Levites, and all they that were come out of the captivity unto Jerusalem; and appointed the Levites, from twenty years old and upward, to set forward the work of the house of the LORD.

eorum ad templum Dei in Hierusalem mense secundo coeperunt Zorobabel filius Salathihel et Iosue filius Iosedech et reliqui de fratribus eorum sacerdotes et Leuitae et omnes qui uenerant de captiuitate in Hierusalem et constituerunt Leui-

3:9 Then stood Jeshua with his sons and his brethren, tas a uiginti annis et supra ut urguerent opus Domini 3:9

Kadmiel and his sons, the sons of Judah, together, to set forward the workmen in the house of God: the sons of

Henadad, with their sons and their brethren the Levites. stetitque Iosue filii eius et fratres eius Cedmihel et filii eius et filii Iuda quasi unus ut instarent super eos qui faciebant opus in templo Dei filii Enadad filii eorum et fratres eorum Leuitae 3:10 fundato igitur a cementariis templo Domini 3:10 And when the builders laid the foundation of the tem-

ple of the LORD, they set the priests in their apparel with trumpets, and the Levites the sons of Asaph with cymbals, to praise the LORD, after the ordinance of David king of Israel.

steterunt sacerdotes in ornatu suo cum tubis et Leuitae filii Asaph in cymbalis ut laudarent Deum per manus Dauid regis Israhel 3:11 et concinebant in hymnis et confessione 3:11 And they sang together by course in praising and giv-

ing thanks unto the LORD; because he is good, for his mercy endureth for ever toward Israel. And all the people shouted with a great shout, when they praised the LORD, because the foundation of the house of the LORD was laid.

Domino quoniam bonus quoniam in aeternum misericordia eius super Israhel omnis quoque populus uociferabatur clamore magno in laudando Dominum eo quod fundatum esset templum Domini 3:12 plurimi etiam de sacerdotibus et Le- 3:12 But many of the priests and Levites and chief of the

fathers, who were ancient men, that had seen the first house, when the foundation of this house was laid before their eyes, wept with a loud voice; and many shouted aloud for joy:

uitis et principes patrum seniores qui uiderant templum prius cum fundatum esset et hoc templum in oculis eorum flebant uoce magna et multi uociferantes in laetitia eleuabant uocem 3:13 nec poterat quisquam agnoscere uocem clamoris 3:13 So that the people could not discern the noise of the

shout of joy from the noise of the weeping of the people: for the people shouted with a loud shout, and the noise was

laetantium et uocem fletus populi commixtim enim populus heard afar off. uociferabatur clamore magno et uox audiebatur procul

4:1 audierunt autem hostes Iudae et Beniamin quia filii ca- 4:1 Now when the adversaries of Judah and Benjamin

heard that the children of the captivity builded the temple unto the LORD God of Israel;

ptiuitatis aedificarent templum Domino Deo Israhel 4:2 et

4:2 Then they came to Zerubbabel, and to the chief of the fathers, and said unto them, Let us build with you: for we seek your God, as ye do; and we do sacrifice unto him since the days of Esarhaddon king of Assur, which brought us up hither.

accedentes ad Zorobabel et ad principes patrum dixerunt eis aedificemus uobiscum quia ita ut uos quaerimus Deum uestrum ecce nos immolamus uictimas ex diebus Asoraddan

regis Assur qui adduxit nos huc 4:3 et dixit eis Zorobabel 4:3 But Zerubbabel, and Jeshua, and the rest of the chief of the fathers of Israel, said unto them, Ye have nothing to do with us to build an house unto our God; but we ourselves together will build unto the LORD God of Israel, as king Cyrus the king of Persia hath commanded us.

et Iosue et reliqui principes patrum Israhel non est uobis et nobis ut aedificemus domum Deo nostro sed nos ipsi soli aedificabimus Domino Deo nostro sicut praecepit nobis rex Cyrus rex Persarum 4:4 factum est igitur ut populus terrae 4:4 Then the people of the land weakened the hands of the

people of Judah, and troubled them in building,

inpediret manus populi Iudae et turbaret eos in aedificando

4:5 conduxerunt quoque aduersum eos consiliatores ut de- 4:5 And hired counsellors against them, to frustrate their purpose, all the days of Cyrus king of Persia, even until the reign of Darius king of Persia.

struerent consilium eorum omnibus diebus Cyri regis Persarum et usque ad regnum Darii regis Persarum 4:6 in regno 4:6 And in the reign of Ahasuerus, in the beginning of

his reign, wrote they unto him an accusation against the inhabitants of Judah and Jerusalem.

autem Asueri principio regni eius scripserunt accusationem aduersum habitatores Iudae et Hierusalem 4:7 et in diebus 4:7 And in the days of Artaxerxes wrote Bishlam, Mithre-

dath, Tabeel, and the rest of their companions, unto Artaxerxes king of Persia; and the writing of the letter was written in the Syrian tongue, and interpreted in the Syrian tongue.

Artarxersis scripsit Beselam Mitridatis et Tabel et reliqui qui erant in consilio eorum ad Artarxersen regem Persarum epistula autem accusationis scripta erat syriace et legebatur ser-

Biblia Sacra Vulgata

4:8 Rehum the chancellor and Shimshai the scribe wrote a mone syro 4:8 Reum Beelteem et Samsai scriba scripserunt

letter against Jerusalem to Artaxerxes the king in this sort:

epistulam unam de Hierusalem Artarxersi regi huiuscemodi

4:9 Then wrote Rehum the chancellor, and Shimshai the 4:9 Reum Beelteem et Samsai scriba et reliqui consiliatores

scribe, and the rest of their companions; the Dinaites, the Apharsathchites, the Tarpelites, the Apharsites, the Archevites, the Babylonians, the Susanchites, the Dehavites, and the Elamites,

eorum Dinei et Apharsathei Terphalei Apharsei Erchuei Ba-

4:10 And the rest of the nations whom the great and noble bylonii Susannechei Deaei Aelamitae 4:10 et ceteri de gen-

Asnapper brought over, and set in the cities of Samaria, and the rest that are on this side the river, and at such a

time. tibus quas transtulit Asennaphar magnus et gloriosus et habi-

tare eas fecit in ciuitatibus Samariae et in reliquis regionibus

4:11 This is the copy of the letter that they sent unto him, trans Flumen in pace 4:11 hoc est exemplar epistulae quam even unto Artaxerxes the king; Thy servants the men on this side the river, and at such a time.

miserunt ad eum Artarxersi regi serui tui uiri qui sunt trans

4:12 Be it known unto the king, that the Jews which came Fluuium salutem dicunt 4:12 notum sit regi quia Iudaei qui

up from thee to us are come unto Jerusalem, building the rebellious and the bad city, and have set up the walls

thereof, and joined the foundations. ascenderunt a te ad nos uenerunt in Hierusalem ciuitatem re-

bellem et pessimam quam aedificant extruentes muros eius

4:13 Be it known now unto the king, that, if this city be et parietes conponentes 4:13 nunc igitur notum sit regi quia

builded, and the walls set up again, then will they not pay toll, tribute, and custom, and so thou shalt endamage the

revenue of the kings. si ciuitas illa aedificata fuerit et muri eius instaurati tributum

et uectigal et annuos rediius non dabunt et usque ad reges

4:14 Now because we have maintenance from the king’s haec noxa perueniet 4:14 nos ergo memores salis quod in

palace, and it was not meet for us to see the king’s dishonour, therefore have we sent and certified the king;

palatio comedimus et quia laesiones regis uidere nefas duci-

4:15 That search may be made in the book of the records mus idcirco misimus et nuntiauimus regi 4:15 ut recenseas

of thy fathers: so shalt thou find in the book of the records, and know that this city is a rebellious city, and hurtful unto kings and provinces, and that they have moved sedition within the same of old time: for which cause was this city destroyed.

in libris historiarum patrum tuorum et inuenies scriptum in commentariis et scies quoniam urbs illa urbs rebellis est et nocens regibus et prouinciis et bella concitant in ea ex die-

4:16 We certify the king that, if this city be builded again, bus antiquis quam ob rem et ciuitas ipsa destructa est 4:16 and the walls thereof set up, by this means thou shalt have no portion on this side the river.

nuntiamus nos regi quoniam si ciuitas illa aedificata fuerit et muri ipsius instaurati possessionem trans Fluuium non ha-

4:17 Then sent the king an answer unto Rehum the chan- bebis 4:17 uerbum misit rex ad Reum Beelteem et Samsai

cellor, and to Shimshai the scribe, and to the rest of their companions that dwell in Samaria, and unto the rest be-

yond the river, Peace, and at such a time. scribam et ad reliquos qui erant in consilio eorum habitatores

Samariae et ceteris trans Fluuium salutem dicens et pacem

4:18 The letter which ye sent unto us hath been plainly 4:18 accusationem quam misistis ad nos manifeste lecta est

read before me.

4:19 And I commanded, and search hath been made, and coram me 4:19 et a me praeceptum est et recensuerunt in-

it is found that this city of old time hath made insurrection against kings, and that rebellion and sedition have been

made therein. ueneruntque quoniam ciuitas illa a diebus antiquis aduersum 4:20 There have been mighty kings also over Jerusalem, reges rebellat et seditiones et proelia concitantur in ea 4:20

which have ruled over all countries beyond the river; and toll, tribute, and custom, was paid unto them.

nam et reges fortissimi fuerunt in Hierusalem qui et dominati sunt omni regioni quae trans Fluuium est tributum quoque et uectigal et rediius accipiebant 4:21 nunc ergo audite sen- 4:21 Give ye now commandment to cause these men to

cease, and that this city be not builded, until another commandment shall be given from me.

tentiam ut prohibeatis uiros illos et urbs illa non aedificetur donec si forte a me iussum fuerit 4:22 uidete ne neglegenter 4:22 Take heed now that ye fail not to do this: why should

damage grow to the hurt of the kings?

hoc impleatis et paulatim crescat malum contra reges 4:23 4:23 Now when the copy of king Artaxerxes’ letter was

read before Rehum, and Shimshai the scribe, and their companions, they went up in haste to Jerusalem unto the

itaque exemplum edicti Artarxersis regis lectum est coram Jews, and made them to cease by force and power. Reum et Samsai scriba et consiliariis eorum et abierunt festini in Hierusalem ad Iudaeos et prohibuerunt eos in brachio et robore 4:24 tunc intermissum est opus domus Dei in Hie- 4:24 Then ceased the work of the house of God which is at

Jerusalem. So it ceased unto the second year of the reign of Darius king of Persia.

rusalem et non fiebat usque ad annum secundum regni Darii regis Persarum

5:1 prophetauerunt autem Aggeus propheta et Zaccharias 5:1 Then the prophets, Haggai the prophet, and Zechariah

the son of Iddo, prophesied unto the Jews that were in Judah and Jerusalem in the name of the God of Israel, even

filius Addo prophetantes ad Iudaeos qui erant in Iudaea et unto them. Hierusalem in nomine Dei Israhel 5:2 tunc surrexerunt Zo- 5:2 Then rose up Zerubbabel the son of Shealtiel, and

Jeshua the son of Jozadak, and began to build the house of God which is at Jerusalem: and with them were the

robabel filius Salathihel et Iosue filius Iosedech et coeperunt prophets of God helping them. aedificare templum Dei in Hierusalem et cum eis prophetae Dei adiuuantes eos 5:3 in ipso tempore uenit ad eos Tatan- 5:3 At the same time came to them Tatnai, governor on this

side the river, and Shetharboznai and their companions, and said thus unto them, Who hath commanded you to

nai qui erat dux trans Flumen et Starbuzannai et consiliarii build this house, and to make up this wall? eorum sicque dixerunt eis quis dedit uobis consilium ut domum hanc aedificaretis et muros hos instauraretis 5:4 ad 5:4 Then said we unto them after this manner, What are

the names of the men that make this building?

quod respondimus eis quae essent nomina hominum auctorum illius aedificationis 5:5 oculus autem Dei eorum fac- 5:5 But the eye of their God was upon the elders of the

Jews, that they could not cause them to cease, till the matter came to Darius: and then they returned answer by letter

tus est super senes Iudaeorum et non potuerunt inhibere eos concerning this matter. placuitque ut res ad Darium referretur et tunc satisfacerent aduersus accusationem illam 5:6 exemplar epistulae quam 5:6 The copy of the letter that Tatnai, governor on this

side the river, and Shetharboznai and his companions the Apharsachites, which were on this side the river, sent unto

misit Tatannai dux regionis trans Flumen et Starbuzannai et Darius the king: consiliatores eius Apharsacei qui erant trans Flumen ad Da-

rium regem 5:7 sermo quem miserant ei sic scriptus erat 5:7 They sent a letter unto him, wherein was written thus;

Unto Darius the king, all peace.

Dario regi pax omnis 5:8 notum sit regi isse nos ad Iu- 5:8 Be it known unto the king, that we went into the

province of Judea, to the house of the great God, which is builded with great stones, and timber is laid in the walls,

daeam prouinciam ad domum Dei magni quae aedificatur and this work goeth fast on, and prospereth in their hands.

lapide inpolito et ligna ponuntur in parietibus opusque illud diligenter extruitur et crescit in manibus eorum 5:9 in- 5:9 Then asked we those elders, and said unto them thus,

Who commanded you to build this house, and to make up these walls?

terrogauimus ergo senes illos et ita diximus eis quis dedit uobis potestatem ut domum hanc aedificaretis et muros in-

Biblia Sacra Vulgata

5:10 We asked their names also, to certify thee, that we stauraretis 5:10 sed et nomina eorum quaesiuimus ab eis ut

might write the names of the men that were the chief of them.

nuntiaremus tibi quae scripsimus nomina uirorum qui sunt

5:11 And thus they returned us answer, saying, We are principes in eis 5:11 huiuscemodi autem sermonem respon-

the servants of the God of heaven and earth, and build the house that was builded these many years ago, which

a great king of Israel builded and set up. derunt nobis dicentes nos sumus serui Dei caeli et terrae et aedificamus templum quod erat extructum ante hos annos multos quodque rex Israhel magnus aedificauerat et extruxe-

5:12 But after that our fathers had provoked the God of rat 5:12 postquam autem ad iracundiam prouocauerunt pa-

heaven unto wrath, he gave them into the hand of Nebuchadnezzar the king of Babylon, the Chaldean, who destroyed this house, and carried the people away into Babylon.

tres nostri Deum caeli et tradidit eos in manu Nabuchodonosor regis Babylonis Chaldei domum quoque hanc destruxit

5:13 But in the first year of Cyrus the king of Babylon the et populum eius transtulit in Babylonem 5:13 anno autem

same king Cyrus made a decree to build this house of God.

primo Cyri regis Babylonis Cyrus rex proposuit edictum ut

5:14 And the vessels also of gold and silver of the house domus Dei aedificaretur 5:14 nam et uasa templi Dei auof God, which Nebuchadnezzar took out of the temple that was in Jerusalem, and brought them into the temple of Babylon, those did Cyrus the king take out of the temple of Babylon, and they were delivered unto one, whose name was Sheshbazzar, whom he had made governor;

rea et argentea quae Nabuchodonosor tulerat de templo quod erat in Hierusalem et asportauerat ea in templum Babylonis protulit Cyrus rex de templo Babylonis et data sunt Sasabas-

5:15 And said unto him, Take these vessels, go, carry them sar uocabulo quem et principem constituit 5:15 dixitque ei

into the temple that is in Jerusalem, and let the house of God be builded in his place.

haec uasa tolle et uade et pone ea in templo quod est in Hie-

5:16 Then came the same Sheshbazzar, and laid the foun- rusalem et domus Dei aedificetur in loco suo 5:16 tunc ita-

dation of the house of God which is in Jerusalem: and since that time even until now hath it been in building, and

yet it is not finished. que Sasabassar ille uenit et posuit fundamenta templi Dei in

Hierusalem et ex eo tempore usque nunc aedificatur et nec-

5:17 Now therefore, if it seem good to the king, let there be dum conpletum est 5:17 nunc ergo si uidetur regi bonum search made in the king’s treasure house, which is there at Babylon, whether it be so, that a decree was made of Cyrus the king to build this house of God at Jerusalem, and let the king send his pleasure to us concerning this matter.

recenseat in bibliotheca regis quae est in Babylone utrumnam a Cyro rege iussum sit ut aedificaretur domus Dei in Hierusalem et uoluntatem regis super hac re mittat ad nos

6:1 Then Darius the king made a decree, and search was 6:1 tunc Darius rex praecepit et recensuerunt in bibliot-

made in the house of the rolls, where the treasures were laid up in Babylon.

heca librorum qui erant repositi in Babylone 6:2 et inuen-

6:2 And there was found at Achmetha, in the palace that is in the province of the Medes, a roll, and therein was a record thus written:

tum est in Ecbathanis quod est castrum in Madena prouincia uolumen unum talisque scriptus erat in eo commenta-

6:3 In the first year of Cyrus the king the same Cyrus rius 6:3 anno primo Cyri regis Cyrus rex decreuit ut domus

the king made a decree concerning the house of God at Jerusalem, Let the house be builded, the place where they offered sacrifices, and let the foundations thereof be strongly laid; the height thereof threescore cubits, and the breadth thereof threescore cubits;

Dei quae est in Hierusalem aedificaretur in loco ubi immolent hostias et ut ponant fundamenta subportantia altitudinem cubitorum sexaginta et latitudinem cubitorum sexaginta

6:4 With three rows of great stones, and a row of new tim- 6:4 ordines de lapidibus inpolitis tres et sic ordines de lig-

ber: and let the expenses be given out of the king’s house: nis nouis sumptus autem de domo regis dabuntur 6:5 sed 6:5 And also let the golden and silver vessels of the house

of God, which Nebuchadnezzar took forth out of the temple which is at Jerusalem, and brought unto Babylon, be restored, and brought again unto the temple which is at Jerusalem, every one to his place, and place them in the house of God.

et uasa templi Dei aurea et argentea quae Nabuchodonosor tulerat de templo Hierusalem et adtulerat ea in Babylonem reddantur et referantur in templo Hierusalem in locum suum quae et posita sunt in templo Dei 6:6 nunc ergo Tatannai 6:6 Now therefore, Tatnai, governor beyond the river,

Shetharboznai, and your companions the Apharsachites, which are beyond the river, be ye far from thence:

dux regionis quae est trans Flumen Starbuzannai et consiliarii uestri Apharsacei qui estis trans Flumen procul recedite ab illis 6:7 et dimittite fieri templum Dei illud a duce Iudae- 6:7 Let the work of this house of God alone; let the gover-

nor of the Jews and the elders of the Jews build this house of God in his place.

orum et a senioribus eorum domum Dei illam aedificent in loco suo 6:8 sed et a me praeceptum est quid oporteat fieri a 6:8 Moreover I make a decree what ye shall do to the el-

ders of these Jews for the building of this house of God: that of the king’s goods, even of the tribute beyond the river, forthwith expenses be given unto these men, that they be not hindered.

presbyteris Iudaeorum illis ut aedificetur domus Dei scilicet ut de arca regis id est de tributis quae dantur de regione trans Flumen studiose sumptus dentur uiris illis ne inpediatur opus 6:9 quod si necesse fuerit et uitulos et agnos et hedos in holo- 6:9 And that which they have need of, both young bul-

locks, and rams, and lambs, for the burnt offerings of the God of heaven, wheat, salt, wine, and oil, according to the appointment of the priests which are at Jerusalem, let it be given them day by day without fail:

caustum Deo caeli frumentum sal uinum et oleum secundum ritum sacerdotum qui sunt in Hierusalem detur eis per dies singulos ne sit in aliquo querimonia 6:10 et offerant obla- 6:10 That they may offer sacrifices of sweet savours unto

the God of heaven, and pray for the life of the king, and of his sons.

tiones Deo caeli orentque pro uita regis et filiorum eius 6:11

6:11 Also I have made a decree, that whosoever shall alter this word, let timber be pulled down from his house, and being set up, let him be hanged thereon; and let his house be made a dunghill for this.

a me ergo positum est decretum ut omnis homo qui hanc mutauerit iussionem tollatur lignum de domo ipsius et erigatur et configatur in eo domus autem eius publicetur 6:12 Deus 6:12 And the God that hath caused his name to dwell

there destroy all kings and people, that shall put to their hand to alter and to destroy this house of God which is at Jerusalem. I Darius have made a decree; let it be done with speed.

autem qui habitare fecit nomen suum ibi dissipet omnia regna et populum qui extenderit manum suam ut repugnet et dissipet domum Dei illam quae est in Hierusalem ego Darius statui decretum quod studiose impleri uolo 6:13 igitur Ta- 6:13 Then Tatnai, governor on this side the river, Shethar-

boznai, and their companions, according to that which Darius the king had sent, so they did speedily.

tannai dux regionis trans Flumen et Starbuzannai et consiliarii eius secundum quod praeceperat Darius rex sic diligenter exsecuti sunt 6:14 seniores autem Iudaeorum aedificabant 6:14 And the elders of the Jews builded, and they pros-

pered through the prophesying of Haggai the prophet and Zechariah the son of Iddo. And they builded, and finished it, according to the commandment of the God of Israel, and according to the commandment of Cyrus, and Darius, and Artaxerxes king of Persia.

et prosperabantur iuxta prophetiam Aggei prophetae et Zacchariae filii Addo et aedificauerunt et construxerunt iubente Deo Israhel et iubente Cyro et Dario et Artarxerse regibus Persarum 6:15 et conpleuerunt domum Dei istam usque ad 6:15 And this house was finished on the third day of the

month Adar, which was in the sixth year of the reign of Darius the king.

diem tertium mensis adar qui est annus sextus regni Darii re-

Biblia Sacra Vulgata

6:16 And the children of Israel, the priests, and the gis 6:16 fecerunt autem filii Israhel sacerdotes et Leuitae et

Levites, and the rest of the children of the captivity, kept the dedication of this house of God with joy.

reliqui filiorum transmigrationis dedicationem domus Dei in

6:17 And offered at the dedication of this house of God an gaudio 6:17 et obtulerunt in dedicationem domus Dei uitu-

hundred bullocks, two hundred rams, four hundred lambs; and for a sin offering for all Israel, twelve he goats, ac-

cording to the number of the tribes of Israel. los centum arietes ducentos agnos quadringentos hircos ca-

prarum pro peccato totius Israhel duodecim iuxta numerum

6:18 And they set the priests in their divisions, and the tribuum Israhel 6:18 et statuerunt sacerdotes in ordinibus

Levites in their courses, for the service of God, which is at Jerusalem; as it is written in the book of Moses.

suis et Leuitas in uicibus suis super opera Dei in Hierusa-

6:19 And the children of the captivity kept the passover lem sicut scriptum est in libro Mosi 6:19 fecerunt autem upon the fourteenth day of the first month.

filii transmigrationis pascha quartadecima die mensis primi

6:20 For the priests and the Levites were purified together, 6:20 purificati enim fuerant sacerdotes et Leuitae quasi unus

all of them were pure, and killed the passover for all the children of the captivity, and for their brethren the priests,

and for themselves. omnes mundi ad immolandum pascha uniuersis filiis trans-6:21 And the children of Israel, which were come again migrationis et fratribus suis sacerdotibus et sibi 6:21 et co-

out of captivity, and all such as had separated themselves unto them from the filthiness of the heathen of the land, to

seek the LORD God of Israel, did eat, mederunt filii Israhel qui reuersi fuerant de transmigratione

et omnis qui se separauerat a coinquinatione gentium ter-

6:22 And kept the feast of unleavened bread seven days rae ad eos ut quaererent Dominum Deum Israhel 6:22 et with joy: for the LORD had made them joyful, and turned the heart of the king of Assyria unto them, to strengthen their hands in the work of the house of God, the God of Israel.

fecerunt sollemnitatem azymorum septem diebus in laetitia quoniam laetificauerat eos Dominus et conuerterat cor regis Assur ad eos ut adiuuaret manus eorum in opere domus Domini Dei Israhel

7:1 Now after these things, in the reign of Artaxerxes king 7:1 post haec autem uerba in regno Artarxersis regis Perof Persia, Ezra the son of Seraiah, the son of Azariah, the son of Hilkiah,

sarum Ezras filius Saraiae filii Azariae filii Helciae 7:2 filii

7:2 The son of Shallum, the son of Zadok, the son of Ahitub,

7:3 The son of Amariah, the son of Azariah, the son of Sellum filii Sadoc filii Achitob 7:3 filii Amariae filii Aza-

Meraioth,

7:4 The son of Zerahiah, the son of Uzzi, the son of Bukki, riae filii Maraioth 7:4 filii Zaraiae filii Ozi filii Bocci 7:5

7:5 The son of Abishua, the son of Phinehas, the son of Eleazar, the son of Aaron the chief priest:

filii Abisue filii Finees filii Eleazar filii Aaron sacerdotis ab

7:6 This Ezra went up from Babylon; and he was a ready initio 7:6 ipse Ezras ascendit de Babylone et ipse scriba ue-

scribe in the law of Moses, which the LORD God of Israel had given: and the king granted him all his request,

according to the hand of the LORD his God upon him. lox in lege Mosi quam dedit Dominus Deus Israhel et dedit

ei rex secundum manum Domini Dei eius super eum om-

7:7 And there went up some of the children of Israel, and nem petitionem eius 7:7 et ascenderunt de filiis Israhel et

of the priests, and the Levites, and the singers, and the porters, and the Nethinims, unto Jerusalem, in the seventh

year of Artaxerxes the king. de filiis sacerdotum et de filiis Leuitarum et de cantoribus

et de ianitoribus et de Nathinneis in Hierusalem anno sep-

7:8 And he came to Jerusalem in the fifth month, which timo Artarxersis regis 7:8 et uenerunt in Hierusalem mense

was in the seventh year of the king.

7:9 For upon the first day of the first month began he to go quinto ipse est annus septimus regis 7:9 quia in primo die

up from Babylon, and on the first day of the fifth month came he to Jerusalem, according to the good hand of his

God upon him. mensis primi coepit ascendere de Babylone et in primo mensis quinti uenit in Hierusalem iuxta manum Dei sui bonam super se 7:10 Ezras enim parauit cor suum ut inuestigaret 7:10 For Ezra had prepared his heart to seek the law of

the LORD, and to do it, and to teach in Israel statutes and judgments.

legem Domini et faceret et doceret in Israhel praeceptum et iudicium 7:11 hoc est autem exemplar epistulae edicti 7:11 Now this is the copy of the letter that the king Artax-

erxes gave unto Ezra the priest, the scribe, even a scribe of the words of the commandments of the LORD, and of his

quod dedit rex Artarxersis Ezrae sacerdoti scribae erudito statutes to Israel. in sermonibus et praeceptis Domini et caerimoniis eius in Israhel 7:12 Artarxersis rex regum Ezrae sacerdoti scribae 7:12 Artaxerxes, king of kings, unto Ezra the priest, a

scribe of the law of the God of heaven, perfect peace, and at such a time.

legis Dei caeli doctissimo salutem 7:13 a me decretum est

7:13 I make a decree, that all they of the people of Israel, and of his priests and Levites, in my realm, which are minded of their own freewill to go up to Jerusalem, go with thee.

ut cuicumque placuerit in regno meo de populo Israhel et de sacerdotibus eius et de Leuitis ire in Hierusalem tecum uadat 7:14 a facie enim regis et septem consiliatorum eius missus 7:14 Forasmuch as thou art sent of the king, and of

his seven counsellors, to enquire concerning Judah and Jerusalem, according to the law of thy God which is in

es ut uisites Iudaeam et Hierusalem in lege Dei tui quae est thine hand;

in manu tua 7:15 et ut feras argentum et aurum quod rex 7:15 And to carry the silver and gold, which the king and his counsellors have freely offered unto the God of Israel, whose habitation is in Jerusalem,

et consiliatores eius sponte obtulerunt Deo Israhel cuius in Hierusalem tabernaculum est 7:16 et omne argentum et au- 7:16 And all the silver and gold that thou canst find in all

the province of Babylon, with the freewill offering of the people, and of the priests, offering willingly for the house

rum quodcumque inueneris in uniuersa prouincia Babylo- of their God which is in Jerusalem: nis et populus offerre uoluerit et de sacerdotibus qui sponte obtulerint domui Dei sui quae est in Hierusalem 7:17 li- 7:17 That thou mayest buy speedily with this money bul-

locks, rams, lambs, with their meat offerings and their drink offerings, and offer them upon the altar of the house

bere accipe et studiose eme de hac pecunia uitulos arietes of your God which is in Jerusalem. agnos et sacrificia et libamina eorum et offer ea super altare templi Dei uestri quod est in Hierusalem 7:18 sed et si quid 7:18 And whatsoever shall seem good to thee, and to thy

brethren, to do with the rest of the silver and the gold, that do after the will of your God.

tibi et fratribus tuis placuerit de reliquo argento et auro ut faciatis iuxta uoluntatem Dei uestri facite 7:19 uasa quo- 7:19 The vessels also that are given thee for the service of

the house of thy God, those deliver thou before the God of Jerusalem.

que quae dantur tibi in ministerium domus Dei tui trade in conspectu Dei Hierusalem 7:20 sed et cetera quibus opus 7:20 And whatsoever more shall be needful for the house

of thy God, which thou shalt have occasion to bestow, bestow it out of the king’s treasure house.

fuerit in domo Dei tui quantumcumque necesse est ut expendas dabis de thesauro et de fisco regis 7:21 et a me ego 7:21 And I, even I Artaxerxes the king, do make a decree

to all the treasurers which are beyond the river, that whatsoever Ezra the priest, the scribe of the law of the God of

Artarxersis rex statui atque decreui omnibus custodibus ar- heaven, shall require of you, it be done speedily, cae publicae qui sunt trans Flumen ut quodcumque petierit a uobis Ezras sacerdos scriba legis Dei caeli absque mora detis 7:22 usque ad argenti talenta centum et usque ad fru- 7:22 Unto an hundred talents of silver, and to an hundred

measures of wheat, and to an hundred baths of wine, and to an hundred baths of oil, and salt without prescribing

menti choros centum et usque ad uini batos centum et usque how much. ad batos olei centum sal uero absque mensura 7:23 omne 7:23 Whatsoever is commanded by the God of heaven, let

it be diligently done for the house of the God of heaven: for why should there be wrath against the realm of the king and his sons?

Biblia Sacra Vulgata

quod ad ritum Dei caeli pertinet tribuatur diligenter in domo Dei caeli ne forte irascatur contra regnum regis et filiorum

7:24 Also we certify you, that touching any of the priests eius 7:24 uobisque notum facimus de uniuersis sacerdoti-

and Levites, singers, porters, Nethinims, or ministers of this house of God, it shall not be lawful to impose toll,

tribute, or custom, upon them. bus et Leuitis cantoribus ianitoribus Nathinneis et ministris

domus Dei huius ut uectigal et tributum et annonas non ha-

7:25 And thou, Ezra, after the wisdom of thy God, that beatis potestatem inponendi super eos 7:25 tu autem Ezras

is in thine hand, set magistrates and judges, which may judge all the people that are beyond the river, all such as know the laws of thy God; and teach ye them that know them not.

secundum sapientiam Dei tui quae est in manu tua constitue iudices et praesides ut iudicent omni populo qui est trans Flumen his uidelicet qui nouerunt legem Dei tui sed et in-

7:26 And whosoever will not do the law of thy God, and peritos docete libere 7:26 et omnis qui non fecerit legem

the law of the king, let judgment be executed speedily upon him, whether it be unto death, or to banishment, or

to confiscation of goods, or to imprisonment. Dei tui et legem regis diligenter iudicium erit de eo siue in

mortem siue in exilium siue in condemnationem substantiae

7:27 Blessed be the LORD God of our fathers, which hath eius uel certe in carcerem 7:27 benedictus Dominus Deus

put such a thing as this in the king’s heart, to beautify the house of the LORD which is in Jerusalem:

patrum nostrorum qui dedit hoc in corde regis ut glorifica-

7:28 And hath extended mercy unto me before the king, ret domum Domini quae est in Hierusalem 7:28 et in me and his counsellors, and before all the king’s mighty princes. And I was strengthened as the hand of the LORD my God was upon me, and I gathered together out of Israel chief men to go up with me.

inclinauit misericordiam coram rege et consiliatoribus eius et uniuersis principibus regis potentibus et ego confortatus manu Domini Dei mei quae erat in me congregaui de Israhel principes qui ascenderent mecum

8:1 These are now the chief of their fathers, and this is the 8:1 hii sunt ergo principes familiarum et genealogia eorum

genealogy of them that went up with me from Babylon, in the reign of Artaxerxes the king.

qui ascenderunt mecum in regno Artarxersis regis de Baby-

8:2 Of the sons of Phinehas; Gershom: of the sons of lone 8:2 de filiis Finees Gersom de filiis Ithamar Danihel

Ithamar; Daniel: of the sons of David; Hattush.

8:3 Of the sons of Shechaniah, of the sons of Pharosh; de filiis Dauid Attus 8:3 de filiis Secheniae et de filiis Pha-

Zechariah: and with him were reckoned by genealogy of the males an hundred and fifty.

ros Zaccharias et cum eo numerati sunt uiri centum quin-

8:4 Of the sons of Pahathmoab; Elihoenai the son of Zer- quaginta 8:4 de filiis Phaethmoab Helioenai filius Zareae

ahiah, and with him two hundred males.

8:5 Of the sons of Shechaniah; the son of Jahaziel, and et cum eo ducenti uiri 8:5 de filiis Secheniae filius Hiezi-

with him three hundred males.

8:6 Of the sons also of Adin; Ebed the son of Jonathan, hel et cum eo trecenti uiri 8:6 de filiis Adden Abeth filius

and with him fifty males. 8:7 And of the sons of Elam; Jeshaiah the son of Athaliah, Ionathan et cum eo quinquaginta uiri 8:7 de filiis Helam and with him seventy males.

8:8 And of the sons of Shephatiah; Zebadiah the son of Isaias filius Athaliae et cum eo septuaginta uiri 8:8 de fi-

Michael, and with him fourscore males.

liis Saphatiae Zebedia filius Michahel et cum eo octoginta

8:9 Of the sons of Joab; Obadiah the son of Jehiel, and uiri 8:9 de filiis Ioab Obedia filius Iehihel et cum eo du-

with him two hundred and eighteen males.

8:10 And of the sons of Shelomith; the son of Josiphiah, centi decem et octo uiri 8:10 de filiis Selomith filius Ios-

and with him an hundred and threescore males. 8:11 And of the sons of Bebai; Zechariah the son of Bebai, phiae et cum eo centum sexaginta uiri 8:11 de filiis Bebai and with him twenty and eight males.

Zaccharias filius Bebai et cum eo uiginti octo uiri 8:12 de 8:12 And of the sons of Azgad; Johanan the son of

Hakkatan, and with him an hundred and ten males.

filiis Ezgad Iohanan filius Eccetan et cum eo centum decem uiri 8:13 de filiis Adonicam qui erant nouissimi et haec no- 8:13 And of the last sons of Adonikam, whose names are

these, Eliphelet, Jeiel, and Shemaiah, and with them threescore males.

mina eorum Helifeleth et Heihel et Samaias et cum eis sexaginta uiri 8:14 de filiis Beggui Vthai et Zacchur et cum eo 8:14 Of the sons also of Bigvai; Uthai, and Zabbud, and

with them seventy males.

septuaginta uiri 8:15 congregaui autem eos ad fluuium qui 8:15 And I gathered them together to the river that runneth to Ahava; and there abode we in tents three days: and I viewed the people, and the priests, and found there none

decurrit ad Ahauua et mansimus ibi diebus tribus quaesiui- of the sons of Levi. que in populo et in sacerdotibus de filiis Leui et non inueni ibi 8:16 itaque misi Heliezer et Arihel et Semeam et Helna- 8:16 Then sent I for Eliezer, for Ariel, for Shemaiah, and

for Elnathan, and for Jarib, and for Elnathan, and for Nathan, and for Zechariah, and for Meshullam, chief men;

than et Iarib et alterum Helnathan et Nathan et Zacchariam also for Joiarib, and for Elnathan, men of understanding. et Mesolam principes et Ioarib et Helnathan sapientes 8:17 8:17 And I sent them with commandment unto Iddo the

chief at the place Casiphia, and I told them what they should say unto Iddo, and to his brethren the Nethinims, at the place Casiphia, that they should bring unto us ministers for the house of our God.

et misi eos ad Heddo qui est primus in Casphiae loco et posui in ore eorum uerba quae loquerentur ad Addom et ad fratres eius Nathinneos in loco Casphiae ut adducerent nobis ministros domus Dei nostri 8:18 et adduxerunt nobis per 8:18 And by the good hand of our God upon us they

brought us a man of understanding, of the sons of Mahli, the son of Levi, the son of Israel; and Sherebiah, with his

manum Dei nostri bonam super nos uirum doctissimum de sons and his brethren, eighteen; filiis Moolli filii Leui filii Israhel et Sarabiam et filios eius

et fratres eius decem et octo 8:19 et Asabiam et cum eo 8:19 And Hashabiah, and with him Jeshaiah of the sons of

Merari, his brethren and their sons, twenty;

Isaiam de filiis Merari fratres eius et filios eius uiginti 8:20 8:20 Also of the Nethinims, whom David and the princes

had appointed for the service of the Levites, two hundred and twenty Nethinims: all of them were expressed

et de Nathinneis quos dederat Dauid et principes ad minis- by name. teria Leuitarum Nathinneos ducentos uiginti omnes hii suis nominibus uocabantur 8:21 et praedicaui ibi ieiunium iuxta 8:21 Then I proclaimed a fast there, at the river of Ahava,

that we might afflict ourselves before our God, to seek of him a right way for us, and for our little ones, and for all

fluuium Ahauua ut adfligeremur coram Domino Deo nostro our substance. et peteremus ab eo uiam rectam nobis et filiis nostris uniuersaeque substantiae nostrae 8:22 erubui enim petere regem 8:22 For I was ashamed to require of the king a band of

soldiers and horsemen to help us against the enemy in the way: because we had spoken unto the king, saying, The hand of our God is upon all them for good that seek him; but his power and his wrath is against all them that forsake him.

auxilium et equites qui defenderent nos ab inimico in uia quia dixeramus regi manus Dei nostri est super omnes qui quaerunt eum in bonitate et imperium eius et fortitudo eius et furor super omnes qui derelinquunt eum 8:23 ieiunaui- 8:23 So we fasted and besought our God for this: and he

was intreated of us.

mus autem et rogauimus Deum nostrum pro hoc et euenit nobis prospere 8:24 et separaui de principibus sacerdotum 8:24 Then I separated twelve of the chief of the priests,

Sherebiah, Hashabiah, and ten of their brethren with them,

duodecim Sarabian Asabian et cum eis de fratribus eorum decem 8:25 adpendique eis argentum et aurum et uasa con- 8:25 And weighed unto them the silver, and the gold, and

the vessels, even the offering of the house of our God, which the king, and his counsellors, and his lords, and all Israel there present, had offered:

Biblia Sacra Vulgata

secrata domus Dei nostri quae obtulerat rex et consiliatores eius et principes eius uniuersusque Israhel eorum qui inuenti

8:26 I even weighed unto their hand six hundred and fifty fuerant 8:26 et adpendi in manibus eorum argenti talenta

talents of silver, and silver vessels an hundred talents, and of gold an hundred talents;

sescenta quinquaginta et uasa argentea centum auri centum

8:27 Also twenty basons of gold, of a thousand drams; and talenta 8:27 et crateras aureos uiginti qui habebant solidos

two vessels of fine copper, precious as gold.

millenos et uasa aeris fulgentis optimi duo pulchra ut aurum

8:28 And I said unto them, Ye are holy unto the LORD; 8:28 et dixi eis uos sancti Domini et uasa sancta et argen-

the vessels are holy also; and the silver and the gold are a freewill offering unto the LORD God of your fathers.

tum et aurum quod sponte oblatum est Domino Deo patrum

8:29 Watch ye, and keep them, until ye weigh them before uestrorum 8:29 uigilate et custodite donec adpendatis co-

the chief of the priests and the Levites, and chief of the fathers of Israel, at Jerusalem, in the chambers of the house

of the LORD. ram principibus sacerdotum et Leuitarum et ducibus fami-

liarum Israhel in Hierusalem et thesaurum domus Domini

8:30 So took the priests and the Levites the weight of 8:30 susceperunt autem sacerdotes et Leuitae pondus argenti

the silver, and the gold, and the vessels, to bring them to Jerusalem unto the house of our God.

et auri et uasorum ut deferrent in Hierusalem in domum Dei

8:31 Then we departed from the river of Ahava on the nostri 8:31 promouimus ergo a flumine Ahauua duodecimo

twelfth day of the first month, to go unto Jerusalem: and the hand of our God was upon us, and he delivered us from the hand of the enemy, and of such as lay in wait by the way.

die mensis primi ut pergeremus Hierusalem et manus Dei nostri fuit super nos et liberauit nos de manu inimici et insi-

8:32 And we came to Jerusalem, and abode there three diatoris in uia 8:32 et uenimus Hierusalem et mansimus ibi

days.

8:33 Now on the fourth day was the silver and the gold and diebus tribus 8:33 die autem quarta adpensum est argentum

the vessels weighed in the house of our God by the hand of Meremoth the son of Uriah the priest; and with him was Eleazar the son of Phinehas; and with them was Jozabad the son of Jeshua, and Noadiah the son of Binnui, Levites;

et aurum et uasa in domo Dei nostri per manum Meremoth filii Vriae sacerdotis et cum eo Eleazar filius Finees cumque eis Iozaded filius Iosue et Noadaia filius Bennoi Leuitae

8:34 By number and by weight of every one: and all the 8:34 iuxta numerum et pondus omnium descriptumque est

weight was written at that time. 8:35 Also the children of those that had been carried away, omne pondus in tempore illo 8:35 sed et qui uenerant de which were come out of the captivity, offered burnt offerings unto the God of Israel, twelve bullocks for all Israel, ninety and six rams, seventy and seven lambs, twelve he goats for a sin offering: all this was a burnt offering unto the LORD.

captiuitate filii transmigrationis obtulerunt holocaustomata Deo Israhel uitulos duodecim pro omni Israhel arietes nonaginta sex agnos septuaginta septem hircos pro peccato duo-

8:36 And they delivered the king’s commissions unto the decim omnia in holocaustum Domino 8:36 dederunt autem king’s lieutenants, and to the governors on this side the river: and they furthered the people, and the house of God.

edicta regis satrapis qui erant de conspectu regis et ducibus trans Flumen et eleuauerunt populum et domum Dei

9:1 Now when these things were done, the princes came 9:1 postquam autem haec conpleta sunt accesserunt ad me

to me, saying, The people of Israel, and the priests, and the Levites, have not separated themselves from the people of the lands, doing according to their abominations, even of the Canaanites, the Hittites, the Perizzites, the Jebusites, the Ammonites, the Moabites, the Egyptians, and the Amorites.

principes dicentes non est separatus populus Israhel et sacerdotes et Leuitae a populis terrarum et de abominationibus eorum Chananei uidelicet et Hetthei et Ferezei et Iebusei et Ammanitarum et Moabitarum et Aegyptiorum et Amorreorum 9:2 tulerunt enim de filiabus eorum sibi et filiis suis et 9:2 For they have taken of their daughters for themselves,

and for their sons: so that the holy seed have mingled themselves with the people of those lands: yea, the hand

commiscuerunt semen sanctum cum populis terrarum ma- of the princes and rulers hath been chief in this trespass.

nus etiam principum et magistratuum fuit in transgressione hac prima 9:3 cumque audissem sermonem istum scidi pal- 9:3 And when I heard this thing, I rent my garment and

my mantle, and plucked off the hair of my head and of my beard, and sat down astonied.

lium meum et tunicam et euelli capillos capitis mei et bar-

bae et sedi maerens 9:4 conuenerunt autem ad me omnes 9:4 Then were assembled unto me every one that trembled at the words of the God of Israel, because of the transgression of those that had been carried away; and I sat astonied

qui timebant uerbum Dei Israhel pro transgressione eorum until the evening sacrifice. qui de captiuitate uenerant et ego sedebam tristis usque ad sacrificium uespertinum 9:5 et in sacrificio uespertino sur- 9:5 And at the evening sacrifice I arose up from my heav-

iness; and having rent my garment and my mantle, I fell upon my knees, and spread out my hands unto the LORD

rexi de adflictione mea et scisso pallio et tunica curuaui ge- my God, nua mea et expandi manus meas ad Dominum Deum meum 9:6 et dixi Deus meus confundor et erubesco leuare Deus 9:6 And said, O my God, I am ashamed and blush to lift up

my face to thee, my God: for our iniquities are increased over our head, and our trespass is grown up unto the heav-

meus faciem meam ad te quoniam iniquitates nostrae mul- ens. tiplicatae sunt super caput et delicta nostra creuerunt usque in caelum 9:7 a diebus patrum nostrorum sed et nos ipsi 9:7 Since the days of our fathers have we been in a great

trespass unto this day; and for our iniquities have we, our kings, and our priests, been delivered into the hand of the kings of the lands, to the sword, to captivity, and to a spoil, and to confusion of face, as it is this day.

peccauimus granditer usque ad diem hanc et in iniquitatibus nostris traditi sumus ipsi et reges nostri et sacerdotes nostri in manum regum terrarum in gladium in captiuitatem in rapinam et in confusionem uultus sicut et die hac 9:8 9:8 And now for a little space grace hath been shewed

from the LORD our God, to leave us a remnant to escape, and to give us a nail in his holy place, that our God may lighten our eyes, and give us a little reviving in our bondage.

et nunc quasi parum et ad momentum facta est deprecatio nostra apud Dominum Deum nostrum ut dimitterentur nobis reliquiae et daretur paxillus in loco sancto eius et inluminaret oculos nostros Deus noster et daret nobis uitam modicam in seruitute nostra 9:9 quia serui sumus et in seruitute nostra 9:9 For we were bondmen; yet our God hath not forsaken

us in our bondage, but hath extended mercy unto us in the sight of the kings of Persia, to give us a reviving, to set up the house of our God, and to repair the desolations thereof, and to give us a wall in Judah and in Jerusalem.

non dereliquit nos Deus noster et inclinauit super nos misericordiam coram rege Persarum ut daret nobis uitam et sublimaret domum Dei nostri et extrueret solitudines eius et daret nobis sepem in Iuda et in Hierusalem 9:10 et nunc quid di- 9:10 And now, O our God, what shall we say after this?

for we have forsaken thy commandments,

cemus Deus noster post haec quia dereliquimus mandata tua 9:11 quae praecepisti in manu seruorum tuorum prophetarum 9:11 Which thou hast commanded by thy servants the

prophets, saying, The land, unto which ye go to possess it, is an unclean land with the filthiness of the people of the lands, with their abominations, which have filled it from one end to another with their uncleanness.

dicens terram ad quam uos ingredimini ut possideatis eam terra inmunda est iuxta inmunditiam populorum ceterarum-

Biblia Sacra Vulgata

que terrarum abominationibus eorum qui repleuerunt eam ab

9:12 Now therefore give not your daughters unto their ore usque ad os in coinquinatione sua 9:12 nunc ergo filias

sons, neither take their daughters unto your sons, nor seek their peace or their wealth for ever: that ye may be strong, and eat the good of the land, and leave it for an inheritance to your children for ever.

uestras ne detis filiis eorum et filias eorum non accipiatis filiis uestris et non quaeratis pacem eorum et prosperitatem eorum usque in aeternum ut confortemini et comedatis quae bona sunt terrae et heredes habeatis filios uestros usque in

9:13 And after all that is come upon us for our evil deeds, saeculum 9:13 et post omnia quae uenerunt super nos in and for our great trespass, seeing that thou our God hast punished us less than our iniquities deserve, and hast given

us such deliverance as this; operibus nostris pessimis et in delicto nostro magno quia tu

Deus noster liberasti nos de iniquitate nostra et dedisti no-

9:14 Should we again break thy commandments, and bis salutem sicut est hodie 9:14 ut non conuerteremur et

join in affinity with the people of these abominations? wouldest not thou be angry with us till thou hadst consumed us, so that there should be no remnant nor escaping?

irrita faceremus mandata tua neque matrimonia iungeremus cum populis abominationum istarum numquid iratus es nobis usque ad consummationem ne dimitteres nobis reliquias

9:15 O LORD God of Israel, thou art righteous: for we et salutem 9:15 Domine Deus Israhel iustus tu quoniam de-

remain yet escaped, as it is this day: behold, we are before thee in our trespasses: for we cannot stand before thee

because of this. relicti sumus qui saluaremur sicut die hac ecce coram te sumus in delicto nostro non enim stari potest coram te super hoc

10:1 Now when Ezra had prayed, and when he had con- 10:1 sic ergo orante Ezra et inplorante eo et flente et ia-

fessed, weeping and casting himself down before the house of God, there assembled unto him out of Israel a very great congregation of men and women and children: for the people wept very sore.

cente ante templum Dei collectus est ad eum de Israhel coetus grandis nimis uirorum et mulierum puerorumque et fleuit

10:2 And Shechaniah the son of Jehiel, one of the sons of populus multo fletu 10:2 et respondit Sechenia filius Iehihel

Elam, answered and said unto Ezra, We have trespassed against our God, and have taken strange wives of the people of the land: yet now there is hope in Israel concerning this thing.

de filiis Helam et dixit Ezrae nos praeuaricati sumus in Deum nostrum et duximus uxores alienigenas de populis terrae

10:3 Now therefore let us make a covenant with our God et nunc si est paenitentia Israhel super hoc 10:3 percutia-

to put away all the wives, and such as are born of them, according to the counsel of my lord, and of those that tremble at the commandment of our God; and let it be done according to the law.

mus foedus cum Deo nostro ut proiciamus uniuersas uxores et eos qui de his nati sunt iuxta uoluntatem Domini et eorum

10:4 Arise; for this matter belongeth unto thee: we also qui timent praeceptum Dei nostri secundum legem fiat 10:4

will be with thee: be of good courage, and do it.

surge tuum est decernere nosque erimus tecum confortare et

10:5 Then arose Ezra, and made the chief priests, the fac 10:5 surrexit ergo Ezras et adiurauit principes sacerdo-

Levites, and all Israel, to swear that they should do according to this word. And they sware.

tum Leuitarum et omnem Israhel ut facerent secundum uer-

10:6 Then Ezra rose up from before the house of God, and bum hoc et iurauerunt 10:6 et surrexit Ezras ante domum

went into the chamber of Johanan the son of Eliashib: and when he came thither, he did eat no bread, nor drink water: for he mourned because of the transgression of them that had been carried away.

Dei et abiit ad cubiculum Iohanan filii Eliasib et ingressus est illuc panem non comedit et aquam non bibit lugebat enim in transgressione eorum qui de captiuitate uenerant 10:7 et 10:7 And they made proclamation throughout Judah and

Jerusalem unto all the children of the captivity, that they should gather themselves together unto Jerusalem;

missa est uox in Iuda et in Hierusalem omnibus filiis transmigrationis ut congregarentur in Hierusalem 10:8 et omnis 10:8 And that whosoever would not come within three

days, according to the counsel of the princes and the elders, all his substance should be forfeited, and himself separated from the congregation of those that had been carried away.

qui non uenerit in tribus diebus iuxta consilium principum et seniorum auferetur uniuersa substantia eius et ipse abicietur de coetu transmigrationis 10:9 conuenerunt igitur om- 10:9 Then all the men of Judah and Benjamin gathered

themselves together unto Jerusalem within three days. It was the ninth month, on the twentieth day of the month; and all the people sat in the street of the house of God, trembling because of this matter, and for the great rain.

nes uiri Iuda et Beniamin in Hierusalem tribus diebus ipse est mensis nonus uicesimo die mensis et sedit omnis populus in platea domus Dei trementes pro peccato et pluuiis 10:10 10:10 And Ezra the priest stood up, and said unto them,

Ye have transgressed, and have taken strange wives, to increase the trespass of Israel.

et surrexit Ezras sacerdos et dixit ad eos uos transgressi estis et duxistis uxores alienigenas ut adderetis super delictum Israhel 10:11 et nunc date confessionem Domino Deo pat- 10:11 Now therefore make confession unto the LORD

God of your fathers, and do his pleasure: and separate yourselves from the people of the land, and from the

rum uestrorum et facite placitum eius et separamini a popu- strange wives. lis terrae et ab uxoribus alienigenis 10:12 et respondit uni- 10:12 Then all the congregation answered and said with a

loud voice, As thou hast said, so must we do.

uersa multitudo dixitque uoce magna iuxta uerbum tuum ad nos sic fiat 10:13 uerumtamen quia populus multus est et 10:13 But the people are many, and it is a time of much

rain, and we are not able to stand without, neither is this a work of one day or two: for we are many that have trans-

tempus pluuiae et non sustinemus stare foris et opus non est gressed in this thing. diei unius uel duorum uehementer quippe peccauimus in sermone isto 10:14 constituantur principes in uniuersa multi- 10:14 Let now our rulers of all the congregation stand,

and let all them which have taken strange wives in our cities come at appointed times, and with them the elders of every city, and the judges thereof, until the fierce wrath of our God for this matter be turned from us.

tudine et omnes in ciuitatibus nostris qui duxerunt uxores alienigenas ueniant in temporibus statutis et cum his seniores per ciuitatem et ciuitatem et iudices eius donec auertatur ira Dei nostri a nobis super peccato hoc 10:15 igitur Ionat- 10:15 Only Jonathan the son of Asahel and Jahaziah the

son of Tikvah were employed about this matter: and Meshullam and Shabbethai the Levite helped them.

han filius Asahel et Iaazia filius Thecuae steterunt super hoc et Mesollam et Sebethai Leuites adiuuerunt eos 10:16 fece- 10:16 And the children of the captivity did so. And Ezra

the priest, with certain chief of the fathers, after the house of their fathers, and all of them by their names, were separated, and sat down in the first day of the tenth month to examine the matter.

runtque sic filii transmigrationis et abierunt Ezras sacerdos et uiri principes familiarum in domum patrum suorum et omnes per nomina sua et sederunt in die primo mensis decimi ut quaererent rem 10:17 et consummati sunt omnes uiri qui 10:17 And they made an end with all the men that had

taken strange wives by the first day of the first month.

duxerant uxores alienigenas usque ad diem primam mensis primi 10:18 et inuenti sunt de filiis sacerdotum qui duxe- 10:18 And among the sons of the priests there were found

that had taken strange wives: namely, of the sons of Jeshua the son of Jozadak, and his brethren; Maaseiah,

rant uxores alienigenas de filiis Iosue filii Iosedech et fratres and Eliezer, and Jarib, and Gedaliah. eius Maasia et Eliezer et Iarib et Godolia 10:19 et dederunt 10:19 And they gave their hands that they would put away

their wives; and being guilty, they offered a ram of the flock for their trespass.

Biblia Sacra Vulgata

manus suas ut eicerent uxores suas et pro delicto suo arie-

10:20 And of the sons of Immer; Hanani, and Zebadiah. tem de ouibus offerrent 10:20 et de filiis Emmer Anani et 10:21 And of the sons of Harim; Maaseiah, and Elijah, Zebedia 10:21 et de filiis Erim Masia et Helia et Semeia et

and Shemaiah, and Jehiel, and Uzziah. 10:22 And of the sons of Pashur; Elioenai, Maaseiah, Ish- Hiehihel et Ozias 10:22 et de filiis Phessur Helioenai Mamael, Nethaneel, Jozabad, and Elasah. 10:23 Also of the Levites; Jozabad, and Shimei, and Kela- asia Ismahel Nathanahel Iozabeth et Elasa 10:23 et de fiiah, (the same is Kelita,) Pethahiah, Judah, and Eliezer.

liis Leuitarum Iozabeth et Semei et Celaia ipse est Calita

10:24 Of the singers also; Eliashib: and of the porters; Phataia Iuda et Eliezer 10:24 et de cantoribus Eliasub et de

Shallum, and Telem, and Uri.

10:25 Moreover of Israel: of the sons of Parosh; Ramiah, ianitoribus Sellum et Telem et Vri 10:25 et ex Israhel de fi-

and Jeziah, and Malchiah, and Miamin, and Eleazar, and Malchijah, and Benaiah.

liis Pharos Remia et Ezia et Melchia et Miamin et Eliezer

10:26 And of the sons of Elam; Mattaniah, Zechariah, and et Melchia et Banea 10:26 et de filiis Helam Mathania Zac-Jehiel, and Abdi, and Jeremoth, and Eliah.

10:27 And of the sons of Zattu; Elioenai, Eliashib, Matta- charias et Hiehil et Abdi et Irimoth et Helia 10:27 et de filiis

niah, and Jeremoth, and Zabad, and Aziza.

Zethua Helioenai Eliasib Mathania et Ierimuth et Zabeth et

10:28 Of the sons also of Bebai; Jehohanan, Hananiah, Aziza 10:28 et de filiis Bebai Iohanan Anania Zabbai At-

Zabbai, and Athlai.

10:29 And of the sons of Bani; Meshullam, Malluch, and halai 10:29 et de filiis Bani Mosollam et Melluch et Adaia

Adaiah, Jashub, and Sheal, and Ramoth.

10:30 And of the sons of Pahathmoab; Adna, and Chelal, Iasub et Saal et Ramoth 10:30 et de filiis Phaethmoab Edna

Benaiah, Maaseiah, Mattaniah, Bezaleel, and Binnui, and Manasseh.

et Chalal Banaias Maasias Mathanias Beselehel et Bennui

10:31 And of the sons of Harim; Eliezer, Ishijah, et Manasse 10:31 et de filiis Erem Eliezer Iesue Melchias

Malchiah, Shemaiah, Shimeon,

10:32 Benjamin, Malluch, and Shemariah. Semeias Symeon 10:32 Beniamin Maloch Samarias 10:33

10:33 Of the sons of Hashum; Mattenai, Mattathah, Zabad, Eliphelet, Jeremai, Manasseh, and Shimei.

de filiis Asom Matthanai Matthetha Zabed Elipheleth Iermai

10:34 Of the sons of Bani; Maadai, Amram, and Uel, Manasse Semei 10:34 de filiis Bani Maaddi Amram et Hu-10:35 Benaiah, Bedeiah, Chelluh, hel 10:35 Baneas et Badaias Cheiliau 10:36 Vannia Mari-

10:36 Vaniah, Meremoth, Eliashib,

10:37 Mattaniah, Mattenai, and Jaasau, muth et Eliasib 10:37 Matthanias Mathanai et Iasi 10:38 et

10:38 And Bani, and Binnui, Shimei,

10:39 And Shelemiah, and Nathan, and Adaiah, Bani et Bennui Semei 10:39 et Salmias et Nathan et Adaias 10:40 Machnadebai, Shashai, Sharai, 10:40 Mechnedabai Sisai Sarai 10:41 Ezrel et Selemau Se-

10:41 Azareel, and Shelemiah, Shemariah,

10:42 Shallum, Amariah, and Joseph. meria 10:42 Sellum Amaria Ioseph 10:43 de filiis Nebu Ia-

10:43 Of the sons of Nebo; Jeiel, Mattithiah, Zabad, Zebina, Jadau, and Joel, Benaiah.

ihel Matthathias Zabed Zabina Ieddu et Iohel Banaia 10:44

10:44 All these had taken strange wives: and some of them had wives by whom they had children.

omnes hii acceperunt uxores alienigenas et fuerunt ex eis mulieres quae pepererant filios

Liber Nehemiae

1:1 uerba Neemiae filii Echliae et factum est in mense cas- 1:1 The words of Nehemiah the son of Hachaliah. And it

came to pass in the month Chisleu, in the twentieth year, as I was in Shushan the palace,

leu anno uicesimo et ego eram in Susis castro 1:2 et uenit

1:2 That Hanani, one of my brethren, came, he and certain men of Judah; and I asked them concerning the Jews that had escaped, which were left of the captivity, and concerning Jerusalem.

Anani unus de fratribus meis ipse et uiri ex Iuda et interrogaui eos de Iudaeis qui remanserant et supererant de captiuitate et de Hierusalem 1:3 et dixerunt mihi qui remanserunt 1:3 And they said unto me, The remnant that are left of

the captivity there in the province are in great affliction and reproach: the wall of Jerusalem also is broken down,

et derelicti sunt de captiuitate ibi in prouincia in adflictione and the gates thereof are burned with fire. magna sunt et in obprobrio et murus Hierusalem dissipatus est et portae eius conbustae sunt igni 1:4 cumque audissem 1:4 And it came to pass, when I heard these words, that I

sat down and wept, and mourned certain days, and fasted, and prayed before the God of heaven,

uerba huiuscemodi sedi et fleui et luxi diebus et ieiunabam et orabam ante faciem Dei caeli 1:5 et dixi quaeso Domine 1:5 And said, I beseech thee, O LORD God of heaven, the

great and terrible God, that keepeth covenant and mercy for them that love him and observe his commandments:

Deus caeli fortis magne atque terribilis qui custodis pactum et misericordiam cum his qui te diligunt et custodiunt mandata tua 1:6 fiat auris tua auscultans et oculi tui aperti ut au- 1:6 Let thine ear now be attentive, and thine eyes open,

that thou mayest hear the prayer of thy servant, which I pray before thee now, day and night, for the children of Israel thy servants, and confess the sins of the children of Israel, which we have sinned against thee: both I and my father’s house have sinned.

dias orationem serui tui quam ego oro coram te hodie nocte et die pro filiis Israhel seruis tuis et confiteor pro peccatis filiorum Israhel quibus peccauerunt tibi et ego et domus patris mei peccauimus 1:7 uanitate seducti sumus et non custodi- 1:7 We have dealt very corruptly against thee, and have

not kept the commandments, nor the statutes, nor the judgments, which thou commandedst thy servant Moses.

uimus mandatum et caerimonias et iudicia quae praecepisti Mosi seruo tuo 1:8 memento uerbi quod mandasti Mosi fa- 1:8 Remember, I beseech thee, the word that thou com-

mandedst thy servant Moses, saying, If ye transgress, I will scatter you abroad among the nations:

mulo tuo dicens cum transgressi fueritis ego dispergam uos in populos 1:9 et si reuertamini ad me et custodiatis man- 1:9 But if ye turn unto me, and keep my commandments,

and do them; though there were of you cast out unto the uttermost part of the heaven, yet will I gather them from thence, and will bring them unto the place that I have chosen to set my name there.

data mea et faciatis ea etiam si abducti fueritis ad extrema caeli inde congregabo uos et inducam in locum quem elegi ut habitaret nomen meum ibi 1:10 et ipsi serui tui et popu- 1:10 Now these are thy servants and thy people, whom

thou hast redeemed by thy great power, and by thy strong hand.

lus tuus quos redeeisti in fortitudine tua magna et in manu tua ualida 1:11 obsecro Domine sit auris tua adtendens ad 1:11 O LORD, I beseech thee, let now thine ear be atten-

tive to the prayer of thy servant, and to the prayer of thy servants, who desire to fear thy name: and prosper, I pray thee, thy servant this day, and grant him mercy in the sight of this man. For I was the king’s cupbearer.

orationem serui tui et ad orationem seruorum tuorum qui uolunt timere nomen tuum et dirige seruum tuum hodie et da ei misericordiam ante uirum hunc ego enim eram pincerna regis

2:1 factum est autem in mense nisan anno uicesimo Ar- 2:1 And it came to pass in the month Nisan, in the twentieth year of Artaxerxes the king, that wine was before him: and I took up the wine, and gave it unto the king. Now I

tarxersis regis et uinum erat ante eum et leuaui uinum et dedi had not been beforetime sad in his presence. regi et non eram quasi languidus ante faciem eius 2:2 dixit- 2:2 Wherefore the king said unto me, Why is thy counte-

nance sad, seeing thou art not sick? this is nothing else but sorrow of heart. Then I was very sore afraid,

Biblia Sacra Vulgata

que mihi rex quare uultus tuus tristis cum te aegrotum non uideam non est hoc frustra sed malum nescio quid in corde

2:3 And said unto the king, Let the king live for ever: tuo est et timui ualde ac nimis 2:3 et dixi regi rex in aeter-

why should not my countenance be sad, when the city, the place of my fathers’ sepulchres, lieth waste, and the gates

thereof are consumed with fire? num uiue quare non maereat uultus meus quia ciuitas domus

sepulchrorum patris mei deserta est et portae eius conbus-

2:4 Then the king said unto me, For what dost thou make tae sunt igni 2:4 et ait mihi rex pro qua re postulas et oraui

request? So I prayed to the God of heaven. 2:5 And I said unto the king, If it please the king, and if thy Deum caeli 2:5 et dixi ad regem si uidetur regi bonum et servant have found favour in thy sight, that thou wouldest send me unto Judah, unto the city of my fathers’ sepul-

chres, that I may build it. si placet seruus tuus ante faciem tuam ut mittas me in Iu-

daeam ad ciuitatem sepulchri patris mei et aedificabo eam

2:6 And the king said unto me, (the queen also sitting by 2:6 dixitque mihi rex et regina quae sedebat iuxta eum usque

him,) For how long shall thy journey be? and when wilt thou return? So it pleased the king to send me; and I set

him a time. ad quod tempus erit iter tuum et quando reuerteris et pla-2:7 Moreover I said unto the king, If it please the king, let cuit ante uultum regis et misit me et constitui ei tempus 2:7

letters be given me to the governors beyond the river, that they may convey me over till I come into Judah;

et dixi regi si regi uidetur bonum epistulas det mihi ad duces regionis trans Flumen ut transducant me donec ueniam

2:8 And a letter unto Asaph the keeper of the king’s forest, in Iudaeam 2:8 et epistulam ad Asaph custodem saltus rethat he may give me timber to make beams for the gates of the palace which appertained to the house, and for the wall of the city, and for the house that I shall enter into. And the king granted me, according to the good hand of my God upon me.

gis ut det mihi ligna et tegere possim portas turris domus et muri ciuitatis et domum quam ingressus fuero et dedit mihi

2:9 Then I came to the governors beyond the river, and rex iuxta manum Dei mei bonam mecum 2:9 et ueni ad dugave them the king’s letters. Now the king had sent captains of the army and horsemen with me.

ces regionis trans Flumen dedique eis epistulas regis miserat

2:10 When Sanballat the Horonite, and Tobiah the servant, autem mecum rex principes militum et equites 2:10 et authe Ammonite, heard of it, it grieved them exceedingly that there was come a man to seek the welfare of the chil-

dren of Israel. dierunt Sanaballat Horonites et Tobias seruus ammanites et

contristati sunt adflictione magna quod uenisset homo qui

2:11 So I came to Jerusalem, and was there three days. quaereret prosperitatem filiorum Israhel 2:11 et ueni Hie-2:12 And I arose in the night, I and some few men with rusalem et eram ibi diebus tribus 2:12 et surrexi nocte ego

me; neither told I any man what my God had put in my heart to do at Jerusalem: neither was there any beast with

me, save the beast that I rode upon. et uiri pauci mecum et non indicaui cuiquam quid Deus de-

disset in corde meo ut facerem in Hierusalem et iumentum

2:13 And I went out by night by the gate of the valley, even non erat mecum nisi animal cui sedebam 2:13 et egressus before the dragon well, and to the dung port, and viewed the walls of Jerusalem, which were broken down, and the

gates thereof were consumed with fire. sum per portam Vallis nocte et ante fontem Draconis et ad

portam Stercoris et considerabam murum Hierusalem dis-

2:14 Then I went on to the gate of the fountain, and to the sipatum et portas eius consumptas igni 2:14 et transiui ad

king’s pool: but there was no place for the beast that was under me to pass.

portam Fontis et ad aquaeductum Regis et non erat locus iu-

2:15 Then went I up in the night by the brook, and viewed mento cui sedebam ut transiret 2:15 et ascendi per torren-

the wall, and turned back, and entered by the gate of the valley, and so returned.

tem nocte et considerabam murum et reuersus ueni ad por-

Liber Nehemiae

tam Vallis et redii 2:16 magistratus autem nesciebant quo 2:16 And the rulers knew not whither I went, or what I did;

neither had I as yet told it to the Jews, nor to the priests, nor to the nobles, nor to the rulers, nor to the rest that did

abissem aut quid ego facerem sed et Iudaeis et sacerdotibus the work. et optimatibus et magistratibus et reliquis qui faciebant opus usque ad id locorum nihil indicaueram 2:17 et dixi eis uos 2:17 Then said I unto them, Ye see the distress that we

are in, how Jerusalem lieth waste, and the gates thereof are burned with fire: come, and let us build up the wall of

nostis adflictionem in qua sumus quia Hierusalem deserta Jerusalem, that we be no more a reproach. est et portae eius consumptae sunt igni uenite et aedificemus muros Hierusalem et non simus ultra obprobrium 2:18 2:18 Then I told them of the hand of my God which was

good upon me; as also the king’s words that he had spoken unto me. And they said, Let us rise up and build. So they

et indicaui eis manum Dei mei quod esset bona mecum et strengthened their hands for this good work. uerba regis quae locutus est mihi et aio surgamus et aedifi-

cemus et confortatae sunt manus eorum in bono 2:19 au- 2:19 But when Sanballat the Horonite, and Tobiah the servant, the Ammonite, and Geshem the Arabian, heard it, they laughed us to scorn, and despised us, and said, What

dierunt autem Sanaballat Horonites et Tobias seruus amma- is this thing that ye do? will ye rebel against the king? nites et Gosem Arabs et subsannauerunt nos et despexerunt dixeruntque quae est haec res quam facitis numquid contra regem uos rebellatis 2:20 et reddidi eis sermonem dixique 2:20 Then answered I them, and said unto them, The God

of heaven, he will prosper us; therefore we his servants will arise and build: but ye have no portion, nor right, nor

ad eos Deus caeli ipse nos iuuat et nos serui eius sumus sur- memorial, in Jerusalem. gamus et aedificemus uobis autem non est pars et iustitia et memoria in Hierusalem

3:1 et surrexit Eliasib sacerdos magnus et fratres eius sa- 3:1 Then Eliashib the high priest rose up with his brethren

the priests, and they builded the sheep gate; they sanctified it, and set up the doors of it; even unto the tower of Meah

cerdotes et aedificauerunt portam Gregis ipsi sanctificaue- they sanctified it, unto the tower of Hananeel. runt eam et statuerunt ualuas eius et usque ad turrem centum cubitorum sanctificauerunt eam usque ad turrem Ananehel 3:2 et iuxta eum aedificauerunt uiri Hiericho et iuxta eum ae- 3:2 And next unto him builded the men of Jericho. And

next to them builded Zaccur the son of Imri.

dificauit Zecchur filius Amri 3:3 portam autem Piscium ae- 3:3 But the fish gate did the sons of Hassenaah build, who

also laid the beams thereof, and set up the doors thereof, the locks thereof, and the bars thereof.

dificauerunt filii Asanaa ipsi texerunt eam et statuerunt ualuas eius et seras et uectes et iuxta eos aedificauit Marimuth filius Vriae filii Accus 3:4 et iuxta eos aedificauit Mosol- 3:4 And next unto them repaired Meremoth the son of Uri-

jah, the son of Koz. And next unto them repaired Meshullam the son of Berechiah, the son of Meshezabeel. And

lam filius Barachiae filii Mesezebel et iuxta eos aedificauit next unto them repaired Zadok the son of Baana. Sadoc filius Baana 3:5 et iuxta eos aedificauerunt Thecueni 3:5 And next unto them the Tekoites repaired; but their

nobles put not their necks to the work of their LORD.

optimates autem eorum non subposuerunt colla sua in opere Domini sui 3:6 et portam Veterem aedificauerunt Ioiada fi- 3:6 Moreover the old gate repaired Jehoiada the son of

Paseah, and Meshullam the son of Besodeiah; they laid the beams thereof, and set up the doors thereof, and the

lius Fasea et Mosollam filius Besodia ipsi texerunt eam et locks thereof, and the bars thereof. statuerunt ualuas eius et seras et uectes 3:7 et iuxta eos ae- 3:7 And next unto them repaired Melatiah the Gibeonite,

and Jadon the Meronothite, the men of Gibeon, and of Mizpah, unto the throne of the governor on this side the

dificauit Meletias Gabaonites et Iadon Meronathites uiri de river.

Biblia Sacra Vulgata

Gabaon et Maspha pro duce qui erat in regione trans Flumen

3:8 Next unto him repaired Uzziel the son of Harhaiah, of 3:8 et iuxta eum aedificauit Ezihel filius Araia aurifex et iuxta

the goldsmiths. Next unto him also repaired Hananiah the son of one of the apothecaries, and they fortified Jerusalem

unto the broad wall. eum aedificauit Anania filius pigmentarii et dimiserunt Hie-3:9 And next unto them repaired Rephaiah the son of Hur, rusalem usque ad murum plateae latioris 3:9 et iuxta eum

the ruler of the half part of Jerusalem.

3:10 And next unto them repaired Jedaiah the son of aedificauit Rafaia filius Ahur princeps uici Hierusalem 3:10

Harumaph, even over against his house. And next unto him repaired Hattush the son of Hashabniah.

et iuxta eos aedificauit Ieiada filius Aromath contra domum

3:11 Malchijah the son of Harim, and Hashub the son of suam et iuxta eum aedificauit Attus filius Asebeniae 3:11

Pahathmoab, repaired the other piece, and the tower of the furnaces.

mediam partem uici aedificauit Melchias filius Erem et Asub

3:12 And next unto him repaired Shallum the son of Halo- filius Phaethmoab et turrem Furnorum 3:12 iuxta eum aedi-

hesh, the ruler of the half part of Jerusalem, he and his daughters.

ficauit Sellum filius Alloes princeps mediae partis uici Hie-

3:13 The valley gate repaired Hanun, and the inhabitants rusalem ipse et filiae eius 3:13 et portam Vallis aedificauit

of Zanoah; they built it, and set up the doors thereof, the locks thereof, and the bars thereof, and a thousand cubits

on the wall unto the dung gate. Anun et habitatores Zanoe ipsi aedificauerunt eam et statue-

runt ualuas eius et seras et uectes et mille cubitos in muro

3:14 But the dung gate repaired Malchiah the son of usque ad portam Sterquilinii 3:14 et portam Sterquilinii ae-

Rechab, the ruler of part of Bethhaccerem; he built it, and set up the doors thereof, the locks thereof, and the bars

thereof. dificauit Melchias filius Rechab princeps uici Bethaccharem

ipse aedificauit eam et statuit ualuas eius et seras et uec-

3:15 But the gate of the fountain repaired Shallun the son tes 3:15 et portam Fontis aedificauit Sellum filius Choloof Colhozeh, the ruler of part of Mizpah; he built it, and covered it, and set up the doors thereof, the locks thereof, and the bars thereof, and the wall of the pool of Siloah by the king’s garden, and unto the stairs that go down from the city of David.

ozai princeps pagi Maspha ipse aedificauit eam et texit et statuit ualuas eius et seras et uectes et muros piscinae Siloae in hortum regis et usque ad gradus qui descendunt de ciuitate

3:16 After him repaired Nehemiah the son of Azbuk, the Dauid 3:16 post eum aedificauit Neemias filius Azboc prin-

ruler of the half part of Bethzur, unto the place over against the sepulchres of David, and to the pool that was made,

and unto the house of the mighty. ceps dimidiae partis uici Bethsur usque contra sepulchra Da-

uid et usque ad piscinam quae grandi opere constructa est et

3:17 After him repaired the Levites, Rehum the son of usque ad domum Fortium 3:17 post eum aedificauerunt Le-

Bani. Next unto him repaired Hashabiah, the ruler of the half part of Keilah, in his part.

uitae Reum filius Benni post eum aedificauit Asebias prin-

3:18 After him repaired their brethren, Bavai the son of ceps dimidiae partis uici Ceilae in uico suo 3:18 post eum

Henadad, the ruler of the half part of Keilah.

aedificauerunt fratres eorum Behui filius Enadad princeps

3:19 And next to him repaired Ezer the son of Jeshua, the dimidiae partis Ceila 3:19 et aedificauit iuxta eum Azer fi-

ruler of Mizpah, another piece over against the going up to the armoury at the turning of the wall.

lius Iosue princeps Maspha mensuram secundam contra as-

3:20 After him Baruch the son of Zabbai earnestly re- censum firmissimi anguli 3:20 post eum in monte aedifi-

paired the other piece, from the turning of the wall unto the door of the house of Eliashib the high priest.

cauit Baruch filius Zacchai mensuram secundam ab angulo

3:21 After him repaired Meremoth the son of Urijah the usque ad portam domus Eliasib sacerdotis magni 3:21 post

son of Koz another piece, from the door of the house of Eliashib even to the end of the house of Eliashib.

eum aedificauit Meremuth filius Vriae filii Accus mensuram

Liber Nehemiae

secundam a porta domus Eliasib donec extenderetur domus Eliasib 3:22 et post eum aedificauerunt sacerdotes uiri de 3:22 And after him repaired the priests, the men of the

plain.

campestribus Iordanis 3:23 post eum aedificauit Beniamin 3:23 After him repaired Benjamin and Hashub over

against their house. After him repaired Azariah the son of Maaseiah the son of Ananiah by his house.

et Asub contra domum suam et post eum aedificauit Azarias filius Maasiae filii Ananiae contra domum suam 3:24 3:24 After him repaired Binnui the son of Henadad an-

other piece, from the house of Azariah unto the turning of the wall, even unto the corner.

post eum aedificauit Bennui filius Enadda mensuram secundam a domo Azariae usque ad flexuram et usque ad angulum 3:25 Falel filius Ozi contra flexuram et turrem quae eminet 3:25 Palal the son of Uzai, over against the turning of the

wall, and the tower which lieth out from the king’s high house, that was by the court of the prison. After him Peda-

de domo regis excelsa id est in atrio carceris post eum Pha- iah the son of Parosh. daia filius Pheros 3:26 Nathinnei autem habitabant in Ofel 3:26 Moreover the Nethinims dwelt in Ophel, unto the

place over against the water gate toward the east, and the tower that lieth out.

usque contra portam Aquarum ad orientem et turrem quae

prominebat 3:27 post eum aedificauerunt Thecueni mensu- 3:27 After them the Tekoites repaired another piece, over against the great tower that lieth out, even unto the wall of Ophel.

ram secundam e regione a turre magna et eminenti usque ad

murum templi 3:28 sursum autem a porta Equorum aedifi- 3:28 From above the horse gate repaired the priests, every

one over against his house.

cauerunt sacerdotes unusquisque contra domum suam 3:29 3:29 After them repaired Zadok the son of Immer over

against his house. After him repaired also Shemaiah the son of Shechaniah, the keeper of the east gate.

post eos aedificauit Seddo filius Emmer contra domum suam et post eum aedificauit Semeia filius Secheniae custos portae orientalis 3:30 post eum aedificauit Anania filius Sele- 3:30 After him repaired Hananiah the son of Shelemiah,

and Hanun the sixth son of Zalaph, another piece. After him repaired Meshullam the son of Berechiah over against his chamber. After him repaired Malchiah the goldsmith’s son unto the place of the Nethinims, and of the merchants, over against the gate Miphkad, and to the going up of the corner.

miae et Anon filius Selo sextus mensuram secundam post eum aedificauit Mosollam filius Barachiae contra gazofilacium suum post eum aedificauit Melchias filius aurificis usque ad domum Nathinneorum et scruta uendentium contra portam Iudicialem et usque ad cenaculum Anguli 3:31 et 3:31 And between the going up of the corner unto the

sheep gate repaired the goldsmiths and the merchants.

inter cenaculum Anguli in porta Gregis aedificauerunt artifices et negotiatores

4:1 factum est autem cum audisset Sanaballat quod aedi- 4:1 But it came to pass, that when Sanballat heard that we

builded the wall, he was wroth, and took great indignation, and mocked the Jews.

ficaremus murum iratus est ualde et motus nimis subsannauit Iudaeos 4:2 et dixit coram fratribus suis et frequentia 4:2 And he spake before his brethren and the army of

Samaria, and said, What do these feeble Jews? will they fortify themselves? will they sacrifice? will they make an end in a day? will they revive the stones out of the heaps of the rubbish which are burned?

Samaritanorum quid Iudaei inbecilli faciunt num dimittent eos gentes num sacrificabunt et conplebunt in una die numquid aedificare poterunt lapides de aceruis pulueris qui conbusti sunt 4:3 sed et Tobias Ammanites proximus eius ait 4:3 Now Tobiah the Ammonite was by him, and he said,

Even that which they build, if a fox go up, he shall even break down their stone wall.

aedificent si ascenderit uulpis transiliet murum eorum lapi-

Biblia Sacra Vulgata

4:4 Hear, O our God; for we are despised: and turn their deum 4:4 audi Deus noster quia facti sumus despectio con-

reproach upon their own head, and give them for a prey in the land of captivity:

uerte obprobrium super caput eorum et da eos in despectio-

4:5 And cover not their iniquity, and let not their sin be nem in terra captiuitatis 4:5 ne operias iniquitatem eorum

blotted out from before thee: for they have provoked thee to anger before the builders.

et peccatum eorum coram facie tua non deleatur quia inri-

4:6 So built we the wall; and all the wall was joined to- serunt aedificantes 4:6 itaque aedificauimus murum et con-

gether unto the half thereof: for the people had a mind to work.

iunximus totum usque ad partem dimidiam et prouocatum

4:7 But it came to pass, that when Sanballat, and Tobiah, est cor populi ad operandum 4:7 factum est autem cum auand the Arabians, and the Ammonites, and the Ashdodites, heard that the walls of Jerusalem were made up, and that the breaches began to be stopped, then they were very wroth,

disset Sanaballat et Tobias et Arabes et Ammanitae et Azotii quod obducta esset cicatrix muri Hierusalem et quod coe-

4:8 And conspired all of them together to come and to fight pissent interrupta concludi irati sunt nimis 4:8 et congre-

against Jerusalem, and to hinder it.

gati omnes pariter ut uenirent et pugnarent contra Hierusa-

4:9 Nevertheless we made our prayer unto our God, and lem et molirentur insidias 4:9 et orauimus Deum nostrum

set a watch against them day and night, because of them.

et posuimus custodes super murum die et nocte contra eos

4:10 And Judah said, The strength of the bearers of bur- 4:10 dixit autem Iudas debilitata est fortitudo portantis et hu-

dens is decayed, and there is much rubbish; so that we are not able to build the wall.

mus nimia est et nos non poterimus aedificare murum 4:11

4:11 And our adversaries said, They shall not know, neither see, till we come in the midst among them, and slay them, and cause the work to cease.

et dixerunt hostes nostri nesciant et ignorent donec ueniamus in medio eorum et interficiamus eos et cessare faciamus

4:12 And it came to pass, that when the Jews which dwelt opus 4:12 factum est autem uenientibus Iudaeis qui habita-

by them came, they said unto us ten times, From all places whence ye shall return unto us they will be upon you.

bant iuxta eos et dicentibus nobis per decem uices ex om-

4:13 Therefore set I in the lower places behind the wall, nibus locis quibus uenerant ad nos 4:13 statui in loco post

and on the higher places, I even set the people after their families with their swords, their spears, and their bows.

murum per circuitum populum in ordine cum gladiis suis et

4:14 And I looked, and rose up, and said unto the nobles, lanceis et arcis 4:14 perspexi atque surrexi et aio ad opti-

and to the rulers, and to the rest of the people, Be not ye afraid of them: remember the LORD, which is great and terrible, and fight for your brethren, your sons, and your daughters, your wives, and your houses.

mates et ad magistratus et ad reliquam partem uulgi nolite timere a facie eorum Domini magni et terribilis mementote et pugnate pro fratribus uestris filiis uestris et filiabus uestris

4:15 And it came to pass, when our enemies heard that it uxoribus uestris et domibus 4:15 factum est autem cum au-

was known unto us, and God had brought their counsel to nought, that we returned all of us to the wall, every one

unto his work. dissent inimici nostri nuntiatum esse nobis dissipauit Deus

consilium eorum et reuersi sumus omnes ad muros unus-

4:16 And it came to pass from that time forth, that the half quisque ad opus suum 4:16 et factum est a die illa media of my servants wrought in the work, and the other half of them held both the spears, the shields, and the bows, and the habergeons; and the rulers were behind all the house of Judah.

pars iuuenum eorum faciebant opus et media parata erat ad bellum et lanceae et scuta et arcus et loricae et principes post

4:17 They which builded on the wall, and they that bare eos in omni domo Iuda 4:17 aedificantium in muro et porburdens, with those that laded, every one with one of his hands wrought in the work, and with the other hand held a

weapon. tantium onera et inponentium una manu sua faciebat opus et

Liber Nehemiae

altera tenebat gladium 4:18 aedificantium enim unusquis- 4:18 For the builders, every one had his sword girded by

his side, and so builded. And he that sounded the trumpet was by me.

que gladio erat accinctus renes et aedificabant et clangebant bucina iuxta me 4:19 et dixi ad optimates et ad magistra- 4:19 And I said unto the nobles, and to the rulers, and to

the rest of the people, The work is great and large, and we are separated upon the wall, one far from another.

tus et ad reliquam partem uulgi opus grande est et latum et

nos separati sumus in muro procul alter ab altero 4:20 in 4:20 In what place therefore ye hear the sound of the trum-

pet, resort ye thither unto us: our God shall fight for us.

loco quocumque audieritis clangorem tubae illuc concurrite ad nos Deus noster pugnabit pro nobis 4:21 et nos ipsi fa- 4:21 So we laboured in the work: and half of them held

the spears from the rising of the morning till the stars appeared.

ciamus opus et media pars nostrum teneat lanceas ab ascensu aurorae donec egrediantur astra 4:22 in tempore quoque illo 4:22 Likewise at the same time said I unto the people, Let

every one with his servant lodge within Jerusalem, that in the night they may be a guard to us, and labour on the day.

dixi populo unusquisque cum puero suo maneat in medio Hierusalem et sint uobis uices per noctem et diem ad operandum 4:23 ego autem et fratres mei et pueri mei et cus- 4:23 So neither I, nor my brethren, nor my servants, nor

the men of the guard which followed me, none of us put off our clothes, saving that every one put them off for wash-

todes qui erant post me non deponebamus uestimenta nostra ing. unusquisque tantum nudabatur ad baptismum

5:1 et factus est clamor populi et uxorum eius magnus 5:1 And there was a great cry of the people and of their

wives against their brethren the Jews.

aduersus fratres suos iudaeos 5:2 et erant qui dicerent fi- 5:2 For there were that said, We, our sons, and our daugh-

ters, are many: therefore we take up corn for them, that we may eat, and live.

lii nostri et filiae nostrae multae sunt nimis accipiamus pro pretio eorum frumentum et comedamus et uiuamus 5:3 et 5:3 Some also there were that said, We have mortgaged

our lands, vineyards, and houses, that we might buy corn, because of the dearth.

erant qui dicerent agros nostros et uineas et domos nostras opponamus et accipiamus frumentum in fame 5:4 et alii di- 5:4 There were also that said, We have borrowed money

for the king’s tribute, and that upon our lands and vineyards.

cebant mutuo sumamus pecunias in tributa regis demusque agros nostros et uineas 5:5 et nunc sicut carnes fratrum no- 5:5 Yet now our flesh is as the flesh of our brethren, our

children as their children: and, lo, we bring into bondage our sons and our daughters to be servants, and some of our daughters are brought unto bondage already: neither is it in our power to redeem them; for other men have our lands and vineyards.

strorum sic carnes nostrae sunt sicut filii eorum ita filii nostri ecce nos subiugamus filios nostros et filias nostras in seruitutem et de filiabus nostris sunt famulae nec habemus unde possint rediii et agros nostros et uineas alii possident 5:6 5:6 And I was very angry when I heard their cry and these

words.

et iratus sum nimis cum audissem clamorem eorum secun-

dum uerba haec 5:7 cogitauitque cor meum mecum et in- 5:7 Then I consulted with myself, and I rebuked the nobles, and the rulers, and said unto them, Ye exact usury, every one of his brother. And I set a great assembly against

crepui optimates et magistratus et dixi eis usurasne singuli them. a fratribus uestris exigatis et congregaui aduersus eos contionem magnam 5:8 et dixi eis nos ut scitis redeeimus fra- 5:8 And I said unto them, We after our ability have re-

deemed our brethren the Jews, which were sold unto the heathen; and will ye even sell your brethren? or shall they be sold unto us? Then held they their peace, and found nothing to answer.

tres nostros iudaeos qui uenditi fuerant gentibus secundum possibilitatem nostram et uos igitur uendite fratres uestros

Biblia Sacra Vulgata

et emimus eos et siluerunt nec inuenerunt quid responde-

5:9 Also I said, It is not good that ye do: ought ye not to rent 5:9 dixique ad eos non est bona res quam facitis quare

walk in the fear of our God because of the reproach of the heathen our enemies?

non in timore Dei nostri ambulatis ne exprobretur nobis a

5:10 I likewise, and my brethren, and my servants, might gentibus inimicis nostris 5:10 et ego et fratres mei et pueri

exact of them money and corn: I pray you, let us leave off this usury.

mei commodauimus plurimis pecuniam et frumentum non repetamus in commune istud aes alienum concedamus quod

5:11 Restore, I pray you, to them, even this day, their debetur nobis 5:11 reddite eis hodie agros suos uineas suas

lands, their vineyards, their oliveyards, and their houses, also the hundredth part of the money, and of the corn, the

wine, and the oil, that ye exact of them. oliueta sua et domos suas quin potius et centesimam pecu-

niae frumenti uini et olei quam exigere soletis ab eis date pro

5:12 Then said they, We will restore them, and will re- illis 5:12 et dixerunt reddimus et ab eis nihil quaerimus sic-

quire nothing of them; so will we do as thou sayest. Then I called the priests, and took an oath of them, that they

should do according to this promise. que faciemus ut loqueris et uocaui sacerdotes et adiuraui eos 5:13 Also I shook my lap, and said, So God shake out ut facerent iuxta quod dixeram 5:13 insuper et sinum meum

every man from his house, and from his labour, that performeth not this promise, even thus be he shaken out, and emptied. And all the congregation said, Amen, and praised the LORD. And the people did according to this promise.

excussi et dixi sic excutiat Deus omnem uirum qui non conpleuerit uerbum istud de domo sua et de laboribus suis sic excutiatur et uacuus fiat et dixit uniuersa multitudo amen et

5:14 Moreover from the time that I was appointed to be laudauerunt Deum fecit ergo populus sicut dictum erat 5:14

their governor in the land of Judah, from the twentieth year even unto the two and thirtieth year of Artaxerxes the king, that is, twelve years, I and my brethren have not eaten the bread of the governor.

a die autem illa qua praeceperat mihi ut essem dux in terra Iuda ab anno uicesimo usque ad annum tricesimum secundum Artarxersis regis per annos duodecim ego et fratres mei

5:15 But the former governors that had been before me annonas quae ducibus debebantur non comedimus 5:15 du-

were chargeable unto the people, and had taken of them bread and wine, beside forty shekels of silver; yea, even their servants bare rule over the people: but so did not I, because of the fear of God.

ces autem primi qui fuerant ante me grauauerunt populum et acceperunt ab eis in pane uino et pecunia cotidie siclos quadraginta sed et ministri eorum depresserant populum ego

5:16 Yea, also I continued in the work of this wall, neither autem non feci ita propter timorem Dei 5:16 quin potius in bought we any land: and all my servants were gathered thither unto the work.

opere muri aedificaui et agrum non emi et omnes pueri mei

5:17 Moreover there were at my table an hundred and fifty congregati ad opus erant 5:17 Iudaei quoque et magistratus

of the Jews and rulers, beside those that came unto us from among the heathen that are about us.

centum quinquaginta uiri et qui ueniebant ad nos de genti-

5:18 Now that which was prepared for me daily was one bus quae in circuitu nostro sunt in mensa mea erant 5:18

ox and six choice sheep; also fowls were prepared for me, and once in ten days store of all sorts of wine: yet for all this required not I the bread of the governor, because the bondage was heavy upon this people.

parabatur autem mihi per dies singulos bos unus arietes sex electi exceptis uolatilibus et inter dies decem uina diuersa et alia multa tribuebam insuper et annonas ducatus mei non

5:19 Think upon me, my God, for good, according to all quaesiui ualde enim erat adtenuatus populus 5:19 memento

that I have done for this people.

mei Deus meus in bonum secundum omnia quae feci populo

Liber Nehemiae

huic

6:1 factum est autem cum audisset Sanaballat et Tobia et 6:1 Now it came to pass when Sanballat, and Tobiah, and

Geshem the Arabian, and the rest of our enemies, heard that I had builded the wall, and that there was no breach left therein; (though at that time I had not set up the doors upon the gates;)

Gosem Arabs et ceteri inimici nostri quod aedificassem ego murum et non esset in ipso residua interruptio usque ad tempus autem illud ualuas non posueram in portis 6:2 miserunt 6:2 That Sanballat and Geshem sent unto me, saying,

Come, let us meet together in some one of the villages in the plain of Ono. But they thought to do me mischief.

Sanaballat et Gosem ad me dicentes ueni et percutiamus foedus pariter in uiculis in campo Ono ipsi autem cogitabant ut facerent mihi malum 6:3 misi ergo ad eos nuntios dicens 6:3 And I sent messengers unto them, saying, I am doing

a great work, so that I cannot come down: why should the work cease, whilst I leave it, and come down to you?

opus grande ego facio et non possum descendere ne forte neglegatur cum uenero et descendero ad uos 6:4 miserunt 6:4 Yet they sent unto me four times after this sort; and I

answered them after the same manner.

autem ad me secundum uerbum hoc per quattuor uices et re-

spondi eis iuxta sermonem priorem 6:5 et misit ad me Sa- 6:5 Then sent Sanballat his servant unto me in like manner

the fifth time with an open letter in his hand;

naballat iuxta uerbum prius quinta uice puerum suum et epi-

stulam habebat in manu scriptam hoc modo 6:6 in gentibus 6:6 Wherein was written, It is reported among the heathen, and Gashmu saith it, that thou and the Jews think to rebel: for which cause thou buildest the wall, that thou mayest

auditum est et Gosem dixit quod tu et Iudaei cogitetis re- be their king, according to these words. bellare et propterea aedifices murum et leuare te uelis super eos regem propter quam causam 6:7 et prophetas posueris 6:7 And thou hast also appointed prophets to preach of

thee at Jerusalem, saying, There is a king in Judah: and now shall it be reported to the king according to these words. Come now therefore, and let us take counsel together.

qui praedicent de te in Hierusalem dicentes rex in Iudaea est auditurus est rex uerba haec idcirco nunc ueni ut ineamus consilium pariter 6:8 et misi ad eos dicens non est factum 6:8 Then I sent unto him, saying, There are no such things

done as thou sayest, but thou feignest them out of thine own heart.

secundum uerba haec quae tu loqueris de corde enim tuo

tu conponis haec 6:9 omnes autem hii terrebant nos cogi- 6:9 For they all made us afraid, saying, Their hands shall be weakened from the work, that it be not done. Now therefore, O God, strengthen my hands.

tantes quod cessarent manus nostrae ab opere et quiescere-

mus quam ob causam magis confortaui manus meas 6:10 et 6:10 Afterward I came unto the house of Shemaiah the son of Delaiah the son of Mehetabeel, who was shut up; and he said, Let us meet together in the house of God, within the temple, and let us shut the doors of the temple: for they will come to slay thee; yea, in the night will they come to slay thee.

ingressus sum domum Samaiae filii Dalaiae filii Metabehel secreto qui ait tractemus nobiscum in domo Dei in medio templi et claudamus portas aedis quia uenturi sunt ut interficiant te et nocte uenturi sunt ad occidendum te 6:11 et dixi 6:11 And I said, Should such a man as I flee? and who is

there, that, being as I am, would go into the temple to save his life? I will not go in.

num quisquam similis mei fugit et quis ut ego ingredietur templum et uiuet non ingrediar 6:12 et intellexi quod De- 6:12 And, lo, I perceived that God had not sent him; but

that he pronounced this prophecy against me: for Tobiah and Sanballat had hired him.

us non misisset eum sed quasi uaticinans locutus esset ad me et Tobia et Sanaballat conduxissent eum 6:13 accepe- 6:13 Therefore was he hired, that I should be afraid, and

do so, and sin, and that they might have matter for an evil report, that they might reproach me.

rat enim pretium ut territus facerem et peccarem et haberent

Biblia Sacra Vulgata

6:14 My God, think thou upon Tobiah and Sanballat ac- malum quod exprobrarent mihi 6:14 memento Domine mei

cording to these their works, and on the prophetess Noadiah, and the rest of the prophets, that would have put me

in fear. pro Tobia et Sanaballat iuxta opera eorum talia sed et No-

adiae prophetae et ceterorum prophetarum qui terrebant me

6:15 So the wall was finished in the twenty and fifth day 6:15 conpletus est autem murus uicesimo quinto die mensis

of the month Elul, in fifty and two days.

6:16 And it came to pass, that when all our enemies heard elul quinquaginta duobus diebus 6:16 factum est ergo cum

thereof, and all the heathen that were about us saw these things, they were much cast down in their own eyes: for

they perceived that this work was wrought of our God. audissent omnes inimici nostri ut timerent uniuersae gentes

quae erant in circuitu nostro et conciderent intra semet ip-

6:17 Moreover in those days the nobles of Judah sent sos et scirent quod a Deo factum esset opus hoc 6:17 sed et

many letters unto Tobiah, and the letters of Tobiah came unto them.

in diebus illis multae optimatium Iudaeorum epistulae mit-

6:18 For there were many in Judah sworn unto him, be- tebantur ad Tobiam et a Tobia ueniebant ad eos 6:18 multi

cause he was the son in law of Shechaniah the son of Arah; and his son Johanan had taken the daughter of Meshullam

the son of Berechiah. enim erant in Iudaea habentes iuramentum eius quia gener

erat Secheniae filii Orei et Iohanan filius eius acceperat fi-

6:19 Also they reported his good deeds before me, and liam Mosollam filii Barachiae 6:19 sed et laudabant eum

uttered my words to him. And Tobiah sent letters to put me in fear.

coram me et uerba mea nuntiabant ei et Tobias mittebat epistulas ut terreret me

7:1 Now it came to pass, when the wall was built, and I 7:1 postquam autem aedificatus est murus et posui ual-

had set up the doors, and the porters and the singers and the Levites were appointed,

uas et recensui ianitores et cantores et Leuitas 7:2 praecepi

7:2 That I gave my brother Hanani, and Hananiah the ruler of the palace, charge over Jerusalem: for he was a faithful man, and feared God above many.

Aneni fratri meo et Ananiae principi domus de Hierusalem ipse enim quasi uir uerax et timens Deum plus ceteris ui-

7:3 And I said unto them, Let not the gates of Jerusalem debatur 7:3 et dixi eis non aperiantur portae Hierusalem

be opened until the sun be hot; and while they stand by, let them shut the doors, and bar them: and appoint watches of the inhabitants of Jerusalem, every one in his watch, and every one to be over against his house.

usque ad calorem solis cumque adhuc adsisterent clausae portae sunt et oppilatae et posui custodes de habitatoribus Hierusalem singulos per uices suas et unumquemque contra

7:4 Now the city was large and great: but the people were domum suam 7:4 ciuitas autem erat lata nimis et grandis et

few therein, and the houses were not builded.

populus paruus in medio eius et non erant domus aedifica-

7:5 And my God put into mine heart to gather together tae 7:5 dedit autem Deus in corde meo et congregaui opti-

the nobles, and the rulers, and the people, that they might be reckoned by genealogy. And I found a register of the genealogy of them which came up at the first, and found written therein,

mates et magistratus et uulgum ut recenserem eos et inueni librum census eorum qui ascenderant primum et inuentum

7:6 These are the children of the province, that went up est scriptum in eo 7:6 isti filii prouinciae qui ascenderunt

out of the captivity, of those that had been carried away, whom Nebuchadnezzar the king of Babylon had carried away, and came again to Jerusalem and to Judah, every one unto his city;

de captiuitate migrantium quos transtulerat Nabuchodonosor rex Babylonis et reuersi sunt in Hierusalem et in Iudaeam

7:7 Who came with Zerubbabel, Jeshua, Nehemiah, unusquisque in ciuitatem suam 7:7 qui uenerunt cum Zo-

Azariah, Raamiah, Nahamani, Mordecai, Bilshan, Mispereth, Bigvai, Nehum, Baanah. The number, I say, of

the men of the people of Israel was this; robabel Hiesuae Neemias Azarias Raamias Naamni Mardo-

Liber Nehemiae

cheus Belsar Mespharath Beggoai Naum Baana numerus uirorum populi Israhel 7:8 filii Pharos duo milia centum se- 7:8 The children of Parosh, two thousand an hundred sev-

enty and two.

ptuaginta duo 7:9 filii Saphatiae trecenti septuaginta duo 7:9 The children of Shephatiah, three hundred seventy and

two.

7:10 filii Area sescenti quinquaginta duo 7:11 filii Phaeth- 7:10 The children of Arah, six hundred fifty and two.

7:11 The children of Pahathmoab, of the children of Jeshua and Joab, two thousand and eight hundred and eighteen.

moab filiorum Hiesuae et Ioab duo milia octingenti decem et octo 7:12 filii Helam mille octingenti quinquaginta quat- 7:12 The children of Elam, a thousand two hundred fifty

and four.

tuor 7:13 filii Zethua octingenti quadraginta quinque 7:14 7:13 The children of Zattu, eight hundred forty and five.

7:14 The children of Zaccai, seven hundred and threescore.

filii Zacchai septingenti sexaginta 7:15 filii Bennui sescenti

7:15 The children of Binnui, six hundred forty and eight.

quadraginta octo 7:16 filii Bebai sescenti uiginti octo 7:17 7:16 The children of Bebai, six hundred twenty and eight.

7:17 The children of Azgad, two thousand three hundred twenty and two.

filii Azgad duo milia trecenti uiginti duo 7:18 filii Adoni-

7:18 The children of Adonikam, six hundred threescore and seven.

cam sescenti sexaginta septem 7:19 filii Baggoaim duo mi-

7:19 The children of Bigvai, two thousand threescore and seven.

lia sexaginta septem 7:20 filii Adin sescenti quinquaginta 7:20 The children of Adin, six hundred fifty and five. quinque 7:21 filii Ater filii Ezechiae nonaginta octo 7:22 7:21 The children of Ater of Hezekiah, ninety and eight.

7:22 The children of Hashum, three hundred twenty and eight.

filii Asem trecenti uiginti octo 7:23 filii Besai trecenti ui-

7:23 The children of Bezai, three hundred twenty and four.

ginti quattuor 7:24 filii Areph centum duodecim 7:25 filii 7:24 The children of Hariph, an hundred and twelve.

7:25 The children of Gibeon, ninety and five.

Gabaon nonaginta quinque 7:26 uiri Bethleem et Netupha _{7:26 The men of Bethlehem and Netophah, an hundred}

fourscore and eight.

centum octoginta octo 7:27 uiri Anathoth centum uiginti 7:27 The men of Anathoth, an hundred twenty and eight. octo 7:28 uiri Bethamoth quadraginta duo 7:29 uiri Caria- 7:28 The men of Bethazmaveth, forty and two.

7:29 The men of Kirjathjearim, Chephirah, and Beeroth, seven hundred forty and three.

thiarim Cephira et Beroth septingenti quadraginta tres 7:30

7:30 The men of Ramah and Gaba, six hundred twenty and one.

uiri Rama et Geba sescenti uiginti unus 7:31 uiri Machmas

7:31 The men of Michmas, an hundred and twenty and two.

centum uiginti duo 7:32 uiri Bethel et Hai centum uiginti 7:32 The men of Bethel and Ai, an hundred twenty and

three.

tres 7:33 uiri Nebo alterius quinquaginta duo 7:34 uiri He- 7:33 The men of the other Nebo, fifty and two.

7:34 The children of the other Elam, a thousand two hundred fifty and four.

lam alterius mille ducenti quinquaginta quattuor 7:35 filii

7:35 The children of Harim, three hundred and twenty.

Arem trecenti uiginti 7:36 filii Hiericho trecenti quadra- 7:36 The children of Jericho, three hundred forty and five. ginta quinque 7:37 filii Lod Adid et Ono septingenti uiginti 7:37 The children of Lod, Hadid, and Ono, seven hundred

twenty and one.

unus 7:38 filii Senaa tria milia nongenti triginta 7:39 sa- 7:38 The children of Senaah, three thousand nine hundred

and thirty.

7:39 The priests: the children of Jedaiah, of the house of Jeshua, nine hundred seventy and three.

cerdotes filii Idaia in domo Iosua nongenti septuaginta tres 7:40 filii Emmer mille quinquaginta duo 7:41 filii Phas- 7:40 The children of Immer, a thousand fifty and two.

7:41 The children of Pashur, a thousand two hundred forty and seven.

sur mille ducenti quadraginta septem 7:42 filii Arem mille

7:42 The children of Harim, a thousand and seventeen.

decem et septem Leuitae 7:43 filii Iosue et Cadmihel fi- 7:43 The Levites: the children of Jeshua, of Kadmiel, and

of the children of Hodevah, seventy and four.

liorum 7:44 Oduia septuaginta quattuor cantores 7:45 fi- 7:44 The singers: the children of Asaph, an hundred forty

and eight.

7:45 The porters: the children of Shallum, the children of Ater, the children of Talmon, the children of Akkub, the children of Hatita, the children of Shobai, an hundred thirty and eight.

lii Asaph centum quadraginta octo 7:46 ianitores filii Sel-

7:46 The Nethinims: the children of Ziha, the children of Hashupha, the children of Tabbaoth,

lum filii Ater filii Telmon filii Accub filii Atita filii Sobai

Biblia Sacra Vulgata

7:47 The children of Keros, the children of Sia, the chil- centum triginta octo 7:47 Nathinnei filii Soa filii Asfa filii

dren of Padon,

7:48 The children of Lebana, the children of Hagaba, the Tebaoth 7:48 filii Ceros filii Siaa filii Fado filii Lebana fi-

children of Shalmai,

7:49 The children of Hanan, the children of Giddel, the lii Agaba filii Selmon 7:49 filii Anan filii Geddel filii Gaer

children of Gahar,

7:50 The children of Reaiah, the children of Rezin, the 7:50 filii Raaia filii Rasim filii Necoda 7:51 filii Gezem filii

children of Nekoda, 7:51 The children of Gazzam, the children of Uzza, the children of Phaseah,

Aza filii Fasea 7:52 filii Besai filii Munim filii Nephusim

7:52 The children of Besai, the children of Meunim, the

children of Nephishesim, 7:53 filii Becbuc filii Acupha filii Arur 7:54 filii Besloth fi-

7:53 The children of Bakbuk, the children of Hakupha, the children of Harhur, 7:54 The children of Bazlith, the children of Mehida, the children of Harsha,

lii Meida filii Arsa 7:55 filii Bercos filii Sisara filii Thema

7:55 The children of Barkos, the children of Sisera, the children of Tamah,

7:56 filii Nesia filii Atipha 7:57 filii seruorum Salomonis fi-

7:56 The children of Neziah, the children of Hatipha. 7:57 The children of Solomon’s servants: the children of Sotai, the children of Sophereth, the children of Perida,

lii Sotai filii Sophereth filii Pherida 7:58 filii Iahala filii De-

7:58 The children of Jaala, the children of Darkon, the children of Giddel,

rcon filii Geddel 7:59 filii Saphatia filii Athil filii Phocereth

7:59 The children of Shephatiah, the children of Hattil, the children of Pochereth of Zebaim, the children of Amon.

qui erat ortus ex Sabaim filio Amon 7:60 omnes Nathin-

7:60 All the Nethinims, and the children of Solomon’s servants, were three hundred ninety and two.

nei et filii seruorum Salomonis trecenti nonaginta duo 7:61

7:61 And these were they which went up also from Telmelah, Telharesha, Cherub, Addon, and Immer: but they could not shew their father’s house, nor their seed, whether they were of Israel.

hii sunt autem qui ascenderunt de Thelmella Thelarsa Cherub Addon et Emmer et non potuerunt indicare domum pa-

7:62 The children of Delaiah, the children of Tobiah, the trum suorum et semen suum utrum ex Israhel essent 7:62 children of Nekoda, six hundred forty and two.

filii Dalaia filii Tobia filii Necoda sescenti quadraginta duo

7:63 And of the priests: the children of Habaiah, the chil- 7:63 et de sacerdotibus filii Abia filii Accos filii Berzellai

dren of Koz, the children of Barzillai, which took one of the daughters of Barzillai the Gileadite to wife, and was

called after their name. qui accepit de filiabus Berzellai Galaditis uxorem et uoca-7:64 These sought their register among those that were tus est nomine eorum 7:64 hii quaesierunt scripturam suam

reckoned by genealogy, but it was not found: therefore were they, as polluted, put from the priesthood.

in censu et non inuenerunt et eiecti sunt de sacerdotio 7:65

7:65 And the Tirshatha said unto them, that they should not eat of the most holy things, till there stood up a priest with Urim and Thummim.

dixitque Athersatha eis ut non manducarent de sanctis sanc-

7:66 The whole congregation together was forty and two torum donec staret sacerdos doctus et eruditus 7:66 omnis

thousand three hundred and threescore,

multitudo quasi unus quadraginta duo milia sescenti sexa-

7:67 Beside their manservants and their maidservants, of ginta 7:67 absque seruis et ancillis eorum qui erant septem

whom there were seven thousand three hundred thirty and seven: and they had two hundred forty and five singing

men and singing women. milia trecenti triginta et septem et inter eos cantores et cant-7:68 Their horses, seven hundred thirty and six: their rices ducentae quadraginta quinque 7:68 7:69 cameli qua-

mules, two hundred forty and five: 7:69 Their camels, four hundred thirty and five: six thousand seven hundred and twenty asses.

dringenti triginta quinque asini sex milia septingenti uiginti

7:70 And some of the chief of the fathers gave unto the 7:70 nonnulli autem de principibus familiarum dederunt in

work. The Tirshatha gave to the treasure a thousand drams of gold, fifty basons, five hundred and thirty priests’ gar-

ments. opus Athersatha dedit in thesaurum auri dragmas mille fialas 7:71 And some of the chief of the fathers gave to the trea- quinquaginta tunicas sacerdotales quingentas triginta 7:71

sure of the work twenty thousand drams of gold, and two thousand and two hundred pound of silver.

et de principibus familiarum dederunt in thesaurum operis auri dragmas uiginti milia et argenti minas duo milia ducentas

Liber Nehemiae

7:72 et quod dedit reliquus populus auri dragmas uiginti milia 7:72 And that which the rest of the people gave was twenty

thousand drams of gold, and two thousand pound of silver, and threescore and seven priests’ garments.

et argenti minas duo milia et tunicas sacerdotales sexaginta septem 7:73 habitauerunt autem sacerdotes et Leuitae et ia- 7:73 So the priests, and the Levites, and the porters, and

the singers, and some of the people, and the Nethinims, and all Israel, dwelt in their cities; and when the seventh

nitores et cantores et reliquum uulgus et Nathinnei et omnis month came, the children of Israel were in their cities. Israhel in ciuitatibus suis

8:1 et uenerat mensis septimus filii autem Israhel erant in 8:1 And all the people gathered themselves together as one

man into the street that was before the water gate; and they spake unto Ezra the scribe to bring the book of the law of

ciuitatibus suis congregatusque est omnis populus quasi uir Moses, which the LORD had commanded to Israel. unus ad plateam quae est ante portam Aquarum et dixerunt Ezrae scribae ut adferret librum legis Mosi quam praecepit Dominus Israheli 8:2 adtulit ergo Ezras sacerdos legem 8:2 And Ezra the priest brought the law before the con-

gregation both of men and women, and all that could hear with understanding, upon the first day of the seventh

coram multitudine uirorum et mulierum cunctisque qui po- month. terant intellegere in die prima mensis septimi 8:3 et legit in 8:3 And he read therein before the street that was before

the water gate from the morning until midday, before the men and the women, and those that could understand; and the ears of all the people were attentive unto the book of the law.

eo aperte in platea quae erat ante portam Aquarum de mane usque ad mediam diem in conspectu uirorum et mulierum et sapientium et aures omnis populi erant erectae ad librum 8:4 stetit autem Ezras scriba super gradum ligneum quem 8:4 And Ezra the scribe stood upon a pulpit of wood,

which they had made for the purpose; and beside him stood Mattithiah, and Shema, and Anaiah, and Urijah, and Hilkiah, and Maaseiah, on his right hand; and on his left hand, Pedaiah, and Mishael, and Malchiah, and Hashum, and Hashbadana, Zechariah, and Meshullam.

fecerat ad loquendum et steterunt iuxta eum Matthathia et Sema et Ania et Vria et Helcia et Maasia ad dextram eius et ad sinistram Phadaia Misahel et Melchia et Asum et Asephdana Zaccharia et Mosollam 8:5 et aperuit Ezras librum co- 8:5 And Ezra opened the book in the sight of all the peo-

ple; (for he was above all the people;) and when he opened it, all the people stood up:

ram omni populo super uniuersum quippe populum eminebat et cum aperuisset eum stetit omnis populus 8:6 et bene- 8:6 And Ezra blessed the LORD, the great God. And all

the people answered, Amen, Amen, with lifting up their hands: and they bowed their heads, and worshipped the

dixit Ezras Domino Deo magno et respondit omnis populus LORD with their faces to the ground. amen amen eleuans manus suas et incuruati sunt et adora-

uerunt Deum proni in terram 8:7 porro Hiesue et Baani et 8:7 Also Jeshua, and Bani, and Sherebiah, Jamin, Akkub, Shabbethai, Hodijah, Maaseiah, Kelita, Azariah, Jozabad, Hanan, Pelaiah, and the Levites, caused the people to un-

Serebia Iamin Accub Septhai Odia Maasia Celita Azarias derstand the law: and the people stood in their place. Iozabed Anam Phalaia Leuitae silentium faciebant in populo ad audiendam legem populus autem stabat in gradu suo 8:8 8:8 So they read in the book in the law of God distinctly,

and gave the sense, and caused them to understand the reading.

et legerunt in libro legis Dei distincte et adposite ad intellegendum et intellexerunt cum legeretur 8:9 dixit autem Ne- 8:9 And Nehemiah, which is the Tirshatha, and Ezra the

priest the scribe, and the Levites that taught the people, said unto all the people, This day is holy unto the LORD your God; mourn not, nor weep. For all the people wept, when they heard the words of the law.

emias ipse est Athersatha et Ezras sacerdos scriba et Leuitae interpretantes uniuerso populo dies sanctificatus est Domino Deo nostro nolite lugere et nolite flere flebat enim omnis po-

Biblia Sacra Vulgata

8:10 Then he said unto them, Go your way, eat the fat, and pulus cum audiret uerba legis 8:10 et dixit eis ite comedite

drink the sweet, and send portions unto them for whom nothing is prepared: for this day is holy unto our LORD: neither be ye sorry; for the joy of the LORD is your strength.

pinguia et bibite mulsum et mittite partes ei qui non praeparauit sibi quia sanctus dies Domini est et nolite contristari

8:11 So the Levites stilled all the people, saying, Hold gaudium enim Domini est fortitudo nostra 8:11 Leuitae au-

your peace, for the day is holy; neither be ye grieved.

tem silentium faciebant in omni populo dicentes tacete quia

8:12 And all the people went their way to eat, and to drink, dies sanctus est et nolite dolere 8:12 abiit itaque omnis po-

and to send portions, and to make great mirth, because they had understood the words that were declared unto

them. pulus ut comederet et biberet et mitteret partes et faceret lae-

titiam magnam quia intellexerant uerba quae docuerat eos

8:13 And on the second day were gathered together the 8:13 et in die secundo congregati sunt principes familiarum

chief of the fathers of all the people, the priests, and the Levites, unto Ezra the scribe, even to understand the words

of the law. uniuersi populi sacerdotes et Leuitae ad Ezram scribam ut 8:14 And they found written in the law which the LORD interpretaretur eis uerba legis 8:14 et inuenerunt scriptum

had commanded by Moses, that the children of Israel should dwell in booths in the feast of the seventh month:

in lege praecepisse Dominum in manu Mosi ut habitent filii

8:15 And that they should publish and proclaim in all their Israhel in tabernaculis in die sollemni mense septimo 8:15

cities, and in Jerusalem, saying, Go forth unto the mount, and fetch olive branches, and pine branches, and myrtle branches, and palm branches, and branches of thick trees, to make booths, as it is written.

et ut praedicent et diuulgent uocem in uniuersis urbibus suis et in Hierusalem dicentes egredimini in montem et adferte frondes oliuae et frondes ligni pulcherrimi frondes myrti et ramos palmarum et frondes ligni nemorosi ut fiant taberna-

8:16 So the people went forth, and brought them, and cula sicut scriptum est 8:16 et egressus est populus et ad-

made themselves booths, every one upon the roof of his house, and in their courts, and in the courts of the house of God, and in the street of the water gate, and in the street of the gate of Ephraim.

tulerunt feceruntque sibi tabernacula unusquisque in domate suo et in atriis suis et in atriis domus Dei et in platea portae

8:17 And all the congregation of them that were come Aquarum et in platea portae Ephraim 8:17 fecit ergo uni-

again out of the captivity made booths, and sat under the booths: for since the days of Jeshua the son of Nun unto that day had not the children of Israel done so. And there was very great gladness.

uersa ecclesia eorum qui rediirant de captiuitate tabernacula et habitauerunt in tabernaculis non enim fecerant a diebus Iosue filii Nun taliter filii Israhel usque ad diem illum et fuit

8:18 Also day by day, from the first day unto the last day, laetitia magna nimis 8:18 legit autem in libro legis Dei per

he read in the book of the law of God. And they kept the feast seven days; and on the eighth day was a solemn

assembly, according unto the manner. dies singulos a die primo usque ad diem nouissimum et fecerunt sollemnitatem septem diebus et in die octauo collectum iuxta ritum

9:1 Now in the twenty and fourth day of this month the 9:1 in die autem uicesimo quarto mensis huius conuene-

children of Israel were assembled with fasting, and with sackclothes, and earth upon them.

runt filii Israhel in ieiunio et in saccis et humus super eos

9:2 And the seed of Israel separated themselves from all 9:2 et separatum est semen filiorum Israhel ab omni filio alie-

strangers, and stood and confessed their sins, and the iniquities of their fathers.

nigena et steterunt et confitebantur peccata sua et iniquitates

9:3 And they stood up in their place, and read in the book patrum suorum 9:3 et consurrexerunt ad standum et lege-

of the law of the LORD their God one fourth part of the day; and another fourth part they confessed, and worshipped the LORD their God.

Liber Nehemiae

runt in uolumine legis Domini Dei sui quater in die et qua-

ter confitebantur et adorabant Dominum Deum suum 9:4 9:4 Then stood up upon the stairs, of the Levites, Jeshua, and Bani, Kadmiel, Shebaniah, Bunni, Sherebiah, Bani, and Chenani, and cried with a loud voice unto the LORD

surrexit autem super gradum Leuitarum Iosue et Bani Ced- their God. mihel Sebnia Bani Sarebias Bani Chanani et inclamauerunt

uoce magna Dominum Deum suum 9:5 et dixerunt Leui- 9:5 Then the Levites, Jeshua, and Kadmiel, Bani, Hashabniah, Sherebiah, Hodijah, Shebaniah, and Pethahiah, said, Stand up and bless the LORD your God for ever and ever: and blessed be thy glorious name, which is exalted above all blessing and praise.

tae Iosue et Cedmihel Bonni Asebia Serebia Odoia Sebna Fataia surgite benedicite Domino Deo uestro ab aeterno usque in aeternum et benedicant nomini gloriae tuae excelso in omni benedictione et laude 9:6 tu ipse Domine solus tu 9:6 Thou, even thou, art LORD alone; thou hast made

heaven, the heaven of heavens, with all their host, the earth, and all things that are therein, the seas, and all that is therein, and thou preservest them all; and the host of heaven worshippeth thee.

fecisti caelum caelum caelorum et omnem exercitum eorum terram et uniuersa quae in ea sunt maria et omnia quae in eis sunt et tu uiuificas omnia haec et exercitus caeli te adorat 9:7 tu ipse Domine Deus qui elegisti Abram et eduxisti eum 9:7 Thou art the LORD the God, who didst choose Abram,

and broughtest him forth out of Ur of the Chaldees, and gavest him the name of Abraham;

de igne Chaldeorum et posuisti nomen eius Abraham 9:8

9:8 And foundest his heart faithful before thee, and madest a covenant with him to give the land of the Canaanites, the Hittites, the Amorites, and the Perizzites, and the Jebusites, and the Girgashites, to give it, I say, to his seed, and hast performed thy words; for thou art righteous:

et inuenisti cor eius fidele coram te et percussisti cum eo foedus ut dares ei terram Chananei Chetthei Amorrei et Ferezei et Iebusei et Gergesei ut dares semini eius et implesti uerba tua quoniam iustus es 9:9 et uidisti adflictionem pa- 9:9 And didst see the affliction of our fathers in Egypt, and

heardest their cry by the Red sea;

trum nostrorum in Aegypto clamoremque eorum audisti super mare Rubrum 9:10 et dedisti signa et portenta in Pharao 9:10 And shewedst signs and wonders upon Pharaoh, and

on all his servants, and on all the people of his land: for thou knewest that they dealt proudly against them. So

et in uniuersis seruis eius et in omni populo terrae illius co- didst thou get thee a name, as it is this day. gnouisti enim quia superbe egerant contra eos et fecisti tibi nomen sicut et in hac die 9:11 et mare diuisisti ante eos et 9:11 And thou didst divide the sea before them, so that

they went through the midst of the sea on the dry land; and their persecutors thou threwest into the deeps, as a stone

transierunt per medium maris in sicca persecutores autem into the mighty waters. eorum proiecisti in profundum quasi lapidem in aquas ualidas 9:12 et in columna nubis ductor eorum fuisti per diem 9:12 Moreover thou leddest them in the day by a cloudy

pillar; and in the night by a pillar of fire, to give them light in the way wherein they should go.

et in columna ignis per noctem ut appareret eis uia per quam ingrediebantur 9:13 ad montem quoque Sinai descendisti 9:13 Thou camest down also upon mount Sinai, and spak-

est with them from heaven, and gavest them right judgments, and true laws, good statutes and commandments:

et locutus es cum eis de caelo et dedisti eis iudicia recta et legem ueritatis caerimonias et praecepta bona 9:14 et sab- 9:14 And madest known unto them thy holy sabbath, and

commandedst them precepts, statutes, and laws, by the hand of Moses thy servant:

batum sanctificatum tuum ostendisti eis et mandata et caerimonias et legem praecepisti eis in manu Mosi serui tui 9:15 9:15 And gavest them bread from heaven for their hunger,

and broughtest forth water for them out of the rock for their thirst, and promisedst them that they should go in to

panem quoque de caelo dedisti eis in fame eorum et aquam possess the land which thou hadst sworn to give them.

Biblia Sacra Vulgata

de petra eduxisti eis sitientibus et dixisti eis ut ingrederentur et possiderent terram super quam leuasti manum tuam ut

9:16 But they and our fathers dealt proudly, and hardened traderes eis 9:16 ipsi uero et patres nostri superbe egerunt

their necks, and hearkened not to thy commandments,

et indurauerunt ceruices suas et non audierunt mandata tua

9:17 And refused to obey, neither were mindful of thy 9:17 et noluerunt audire et non sunt recordati mirabilium tuo-

wonders that thou didst among them; but hardened their necks, and in their rebellion appointed a captain to return to their bondage: but thou art a God ready to pardon, gracious and merciful, slow to anger, and of great kindness, and forsookest them not.

rum quae feceras eis et indurauerunt ceruices suas et dederunt caput ut conuerterentur ad seruitutem suam quasi per contentionem tu autem Deus propitius clemens et misericors

9:18 Yea, when they had made them a molten calf, and longanimis et multae miserationis non dereliquisti eos 9:18

said, This is thy God that brought thee up out of Egypt, and had wrought great provocations;

et quidem cum fecissent sibi uitulum conflatilem et dixissent iste est Deus tuus qui eduxit te de Aegypto feceruntque blas-

9:19 Yet thou in thy manifold mercies forsookest them not phemias magnas 9:19 tu autem in misericordiis tuis multis

in the wilderness: the pillar of the cloud departed not from them by day, to lead them in the way; neither the pillar of fire by night, to shew them light, and the way wherein they should go.

non dimisisti eos in deserto columna nubis non recessit ab eis per diem ut duceret eos in uia et columna ignis in nocte

9:20 Thou gavest also thy good spirit to instruct them, and ut ostenderet eis iter per quod ingrederentur 9:20 et spiri-

withheldest not thy manna from their mouth, and gavest them water for their thirst.

tum tuum bonum dedisti qui doceret eos et manna tuum non

9:21 Yea, forty years didst thou sustain them in the wilder- prohibuisti ab ore eorum et aquam dedisti eis in siti 9:21 ness, so that they lacked nothing; their clothes waxed not old, and their feet swelled not.

quadraginta annis pauisti eos in deserto nihilque eis defuit uestimenta eorum non inueterauerunt et pedes eorum non

9:22 Moreover thou gavest them kingdoms and nations, sunt adtriti 9:22 et dedisti eis regna et populos et partitus es

and didst divide them into corners: so they possessed the land of Sihon, and the land of the king of Heshbon, and

the land of Og king of Bashan. eis sortes et possederunt terram Seon et terram regis Esebon 9:23 Their children also multipliedst thou as the stars et terram Og regis Basan 9:23 et filios eorum multiplicasti

of heaven, and broughtest them into the land, concerning which thou hadst promised to their fathers, that they

should go in to possess it. sicut stellas caeli et adduxisti eos ad terram de qua dixeras 9:24 So the children went in and possessed the land, and patribus eorum ut ingrederentur et possiderent 9:24 et ue-

thou subduedst before them the inhabitants of the land, the Canaanites, and gavest them into their hands, with their kings, and the people of the land, that they might do with them as they would.

nerunt filii et possederunt terram et humiliasti coram eis habitatores terrae Chananeos et dedisti eos in manu eorum et reges eorum et populos terrae ut facerent eis sicut placebat

9:25 And they took strong cities, and a fat land, and pos- illis 9:25 ceperunt itaque urbes munitas et humum pinguem

sessed houses full of all goods, wells digged, vineyards, and oliveyards, and fruit trees in abundance: so they did eat, and were filled, and became fat, and delighted themselves in thy great goodness.

et possederunt domos plenas cunctis bonis cisternas ab aliis fabricatas uineas et oliueta et ligna pomifera multa et comederunt et saturati sunt et inpinguati sunt et abundauere

9:26 Nevertheless they were disobedient, and rebelled deliciis in bonitate tua magna 9:26 prouocauerunt autem te

against thee, and cast thy law behind their backs, and slew thy prophets which testified against them to turn them to

thee, and they wrought great provocations. ad iracundiam et recesserunt a te et proiecerunt legem tuam

Liber Nehemiae

post terga sua et prophetas tuos occiderunt qui contestabantur eos ut reuerterentur ad te feceruntque blasphemias gran-

des 9:27 et dedisti eos in manu hostium suorum et adflixe- 9:27 Therefore thou deliveredst them into the hand of their enemies, who vexed them: and in the time of their trouble, when they cried unto thee, thou heardest them from heaven; and according to thy manifold mercies thou gavest them saviours, who saved them out of the hand of their enemies.

runt eos et in tempore tribulationis suae clamauerunt ad te et tu de caelo audisti et secundum miserationes tuas multas dedisti eis saluatores qui saluauerunt eos de manu hostium

suorum 9:28 cumque requieuissent reuersi sunt ut facerent 9:28 But after they had rest, they did evil again before thee: therefore leftest thou them in the land of their enemies, so that they had the dominion over them: yet when they returned, and cried unto thee, thou heardest them from heaven; and many times didst thou deliver them according to thy mercies;

malum in conspectu tuo et dereliquisti eos in manu inimicorum suorum et possederunt eos conuersique sunt et clamauerunt ad te tu autem de caelo audisti et liberasti eos in misericordiis tuis multis temporibus 9:29 et contestatus es 9:29 And testifiedst against them, that thou mightest bring

them again unto thy law: yet they dealt proudly, and hearkened not unto thy commandments, but sinned against thy judgments, (which if a man do, he shall live in them;) and withdrew the shoulder, and hardened their neck, and would not hear.

eos ut reuerterentur ad legem tuam ipsi uero superbe egerunt et non audierunt mandata tua et in iudiciis tuis peccauerunt quae faciet homo et uiuet in eis et dederunt umerum recedentem et ceruicem suam indurauerunt nec audierunt 9:30 9:30 Yet many years didst thou forbear them, and testi-

fiedst against them by thy spirit in thy prophets: yet would they not give ear: therefore gavest thou them into the hand

et protraxisti super eos annos multos et contestatus es eos in of the people of the lands. spiritu tuo per manum prophetarum tuorum et non audierunt et tradidisti eos in manu populorum terrarum 9:31 in mise- 9:31 Nevertheless for thy great mercies’ sake thou didst

not utterly consume them, nor forsake them; for thou art a gracious and merciful God.

ricordiis autem tuis plurimis non fecisti eos in consumptione nec dereliquisti eos quoniam Deus miserationum et clemens tu es 9:32 nunc itaque Deus noster Deus magne fortis et ter- 9:32 Now therefore, our God, the great, the mighty, and

the terrible God, who keepest covenant and mercy, let not all the trouble seem little before thee, that hath come upon us, on our kings, on our princes, and on our priests, and on our prophets, and on our fathers, and on all thy people, since the time of the kings of Assyria unto this day.

ribilis custodiens pactum et misericordiam ne auertas a facie tua omnem laborem qui inuenit nos reges nostros principes nostros et sacerdotes nostros prophetas nostros et patres nostros et omnem populum tuum a diebus regis Assur usque in diem hanc 9:33 et tu iustus in omnibus quae uenerunt super 9:33 Howbeit thou art just in all that is brought upon us;

for thou hast done right, but we have done wickedly:

nos quia ueritatem fecisti nos autem impie egimus 9:34 re- 9:34 Neither have our kings, our princes, our priests, nor

our fathers, kept thy law, nor hearkened unto thy commandments and thy testimonies, wherewith thou didst tes-

ges nostri principes nostri sacerdotes nostri et patres nostri tify against them. non fecerunt legem tuam et non adtenderunt mandata tua et testimonia tua quae testificatus es in eis 9:35 et ipsi in re- 9:35 For they have not served thee in their kingdom, and in

thy great goodness that thou gavest them, and in the large and fat land which thou gavest before them, neither turned

gnis suis bonis et in bonitate tua multa quam dederas eis et they from their wicked works. in terra latissima et pingui quam tradideras in conspectu eorum non seruierunt tibi nec reuersi sunt ab studiis suis pes-

9:36 Behold, we are servants this day, and for the land that simis 9:36 ecce nos ipsi hodie serui sumus et terram quam thou gavest unto our fathers to eat the fruit thereof and the good thereof, behold, we are servants in it:

dedisti patribus nostris ut comederent panem eius et quae

9:37 And it yieldeth much increase unto the kings whom bona sunt eius et nos ipsi serui sumus in ea 9:37 et fruges

thou hast set over us because of our sins: also they have dominion over our bodies, and over our cattle, at their

pleasure, and we are in great distress. eius multiplicantur regibus quos posuisti super nos propter peccata nostra et in corporibus nostris dominantur et in iumentis nostris secundum uoluntatem suam et in tribulatione

9:38 And because of all this we make a sure covenant, and magna sumus 9:38 super omnibus ergo his nos ipsi percuwrite it; and our princes, Levites, and priests, seal unto it.

timus foedus et scribimus et signant principes nostri Leuitae nostri et sacerdotes nostri

10:1 Now those that sealed were, Nehemiah, the Tirshatha, 10:1 signatores autem fuerunt Neemias Athersatha filius the son of Hachaliah, and Zidkijah,

10:2 Seraiah, Azariah, Jeremiah, Achelai et Sedecias 10:2 Saraias Azarias Hieremias 10:3

10:3 Pashur, Amariah, Malchijah,

10:4 Hattush, Shebaniah, Malluch, Phessur Amaria Melchia 10:4 Attus Sebenia Melluc 10:5

10:5 Harim, Meremoth, Obadiah,

10:6 Daniel, Ginnethon, Baruch, Arem Mermuth Obdias 10:6 Danihel Genton Baruch 10:7

10:7 Meshullam, Abijah, Mijamin,

Mosollam Abia Miamin 10:8 Mazia Belga Semaia hii sa-

10:8 Maaziah, Bilgai, Shemaiah: these were the priests.

10:9 And the Levites: both Jeshua the son of Azaniah, cerdotes 10:9 porro Leuitae Iosue filius Azaniae Bennui

Binnui of the sons of Henadad, Kadmiel;

10:10 And their brethren, Shebaniah, Hodijah, Kelita, de filiis Enadad Cedmihel 10:10 et fratres eorum Seche-Pelaiah, Hanan,

10:11 Micha, Rehob, Hashabiah, nia Odeuia Celita Phalaia Anan 10:11 Micha Roob Ase-10:12 Zaccur, Sherebiah, Shebaniah, bia 10:12 Zacchur Serebia Sabania 10:13 Odia Bani Ba-

10:13 Hodijah, Bani, Beninu.

10:14 The chief of the people; Parosh, Pahathmoab, Elam, ninu 10:14 capita populi Pheros Phaethmoab Helam Zethu Zatthu, Bani,

10:15 Bunni, Azgad, Bebai, Bani 10:15 Bonni Azgad Bebai 10:16 Adonia Beggoai

10:16 Adonijah, Bigvai, Adin,

10:17 Ater, Hizkijah, Azzur, Adin 10:17 Ater Ezechia Azur 10:18 Odeuia Asum Be-

10:18 Hodijah, Hashum, Bezai,

sai 10:19 Ares Anathoth Nebai 10:20 Mecphia Mosollam

10:19 Hariph, Anathoth, Nebai,

10:20 Magpiash, Meshullam, Hezir,

Azir 10:21 Mesizabel Sadoc Ieddua 10:22 Felthia Anan

10:21 Meshezabeel, Zadok, Jaddua,

10:22 Pelatiah, Hanan, Anaiah,

Ania 10:23 Osee Anania Asub 10:24 Aloes Phaleam So-

10:23 Hoshea, Hananiah, Hashub,

10:24 Hallohesh, Pileha, Shobek,

bec 10:25 Reum Asebna Madsia 10:26 et Haia Hanam

10:25 Rehum, Hashabnah, Maaseiah,

10:26 And Ahijah, Hanan, Anan,

Anan 10:27 Melluc Arem Baana 10:28 et reliqui de populo

10:27 Malluch, Harim, Baanah.

10:28 And the rest of the people, the priests, the Levites, the porters, the singers, the Nethinims, and all they that had separated themselves from the people of the lands unto the law of God, their wives, their sons, and their daughters, every one having knowledge, and having understanding;

sacerdotes Leuitae ianitores et cantores Nathinnei et omnes qui se separauerunt de populis terrarum ad legem Dei uxo-

10:29 They clave to their brethren, their nobles, and en- res eorum filii eorum et filiae eorum 10:29 omnis qui potered into a curse, and into an oath, to walk in God’s law, which was given by Moses the servant of God, and to observe and do all the commandments of the LORD our Lord, and his judgments and his statutes;

terat sapere spondentes pro fratribus suis optimates eorum et qui ueniebant ad pollicendum et iurandum ut ambularent in lege Dei quam dederat in manu Mosi serui Dei ut facerent et custodirent uniuersa mandata Domini Dei nostri et

Liber Nehemiae

iudicia eius et caerimonias eius 10:30 et ut non daremus fi- 10:30 And that we would not give our daughters unto the

people of the land, not take their daughters for our sons:

lias nostras populo terrae et filias eorum non acciperemus filiis nostris 10:31 populi quoque terrae qui inportant uena- 10:31 And if the people of the land bring ware or any vict-

uals on the sabbath day to sell, that we would not buy it of them on the sabbath, or on the holy day: and that we would leave the seventh year, and the exaction of every debt.

lia et omnia ad usum per diem sabbati ut uendant non accipiemus ab eis in sabbato et in die sanctificata et dimittemus

annum septimum et exactionem uniuersae manus 10:32 et 10:32 Also we made ordinances for us, to charge ourselves yearly with the third part of a shekel for the service of the house of our God;

statuemus super nos praecepta ut demus tertiam partem sicli

per annum ad opus domus Dei nostri 10:33 ad panes pro- 10:33 For the shewbread, and for the continual meat offering, and for the continual burnt offering, of the sabbaths, of the new moons, for the set feasts, and for the holy things, and for the sin offerings to make an atonement for Israel, and for all the work of the house of our God.

positionis et ad sacrificium sempiternum et in holocaustum sempiternum in sabbatis in kalendis in sollemnitatibus et in sanctificatis et pro peccato ut exoretur pro Israhel et in om-

nem usum domus Dei nostri 10:34 sortes ergo misimus su- 10:34 And we cast the lots among the priests, the Levites, and the people, for the wood offering, to bring it into the house of our God, after the houses of our fathers, at times appointed year by year, to burn upon the altar of the LORD our God, as it is written in the law:

per oblatione lignorum inter sacerdotes et Leuitas et populos ut inferrentur in domum Dei nostri per domos patrum nostrorum per tempora a temporibus anni usque ad annum ut arderent super altare Domini Dei nostri sicut scriptum est in lege Mosi 10:35 et ut adferremus primogenita terrae no- 10:35 And to bring the firstfruits of our ground, and the

firstfruits of all fruit of all trees, year by year, unto the house of the LORD:

strae et primitiua uniuersi fructus omnis ligni ab anno in annum in domo Domini 10:36 et primitiua filiorum nostrorum 10:36 Also the firstborn of our sons, and of our cattle, as

it is written in the law, and the firstlings of our herds and of our flocks, to bring to the house of our God, unto the

et pecorum nostrorum sicut scriptum est in lege et primi- priests that minister in the house of our God: tiua boum nostrorum et ouium nostrarum ut offerrentur in domo Dei nostri sacerdotibus qui ministrant in domo Dei nostri 10:37 et primitias ciborum nostrorum et libaminum 10:37 And that we should bring the firstfruits of our

dough, and our offerings, and the fruit of all manner of trees, of wine and of oil, unto the priests, to the chambers of the house of our God; and the tithes of our ground unto the Levites, that the same Levites might have the tithes in all the cities of our tillage.

nostrorum et poma omnis ligni uindemiae quoque et olei adferemus sacerdotibus ad gazofilacium Dei nostri et decimam partem terrae nostrae Leuitis ipsi Leuitae decimas accipient ex omnibus ciuitatibus operum nostrorum 10:38 erit autem 10:38 And the priest the son of Aaron shall be with the

Levites, when the Levites take tithes: and the Levites shall bring up the tithe of the tithes unto the house of our God,

sacerdos filius Aaron cum Leuitis in decimis Leuitarum et to the chambers, into the treasure house. Leuitae offerent decimam partem decimae suae in domum Dei nostri ad gazofilacium in domo thesauri 10:39 ad gazo- 10:39 For the children of Israel and the children of Levi

shall bring the offering of the corn, of the new wine, and the oil, unto the chambers, where are the vessels of the sanctuary, and the priests that minister, and the porters, and the singers: and we will not forsake the house of our God.

filacium enim deportabunt filii Israhel et filii Leui primitias frumenti uini et olei et ibi erunt uasa sanctificata et sacerdotes et cantores et ianitores et ministri et non dimittemus

Biblia Sacra Vulgata

domum Dei nostri

11:1 And the rulers of the people dwelt at Jerusalem: the 11:1 habitauerunt autem principes populi in Hierusalem

rest of the people also cast lots, to bring one of ten to dwell in Jerusalem the holy city, and nine parts to dwell in other

cities. reliqua uero plebs misit sortem ut tollerent unam partem de

decem qui habitaturi essent in Hierusalem in ciuitate sancta

11:2 And the people blessed all the men, that willingly nouem uero partes in ciuitatibus 11:2 benedixit autem po-

offered themselves to dwell at Jerusalem.

pulus omnibus uiris qui se sponte obtulerunt ut habitarent

11:3 Now these are the chief of the province that dwelt in in Hierusalem 11:3 hii sunt itaque principes prouinciae qui

Jerusalem: but in the cities of Judah dwelt every one in his possession in their cities, to wit, Israel, the priests, and the Levites, and the Nethinims, and the children of Solomon’s servants.

habitauerunt in Hierusalem et in ciuitatibus Iuda habitauit unusquisque in possessione sua in urbibus suis Israhel sacer-

11:4 And at Jerusalem dwelt certain of the children of Ju- dotes Leuitae Nathinnei et filii seruorum Salomonis 11:4 et

dah, and of the children of Benjamin. Of the children of Judah; Athaiah the son of Uzziah, the son of Zechariah, the son of Amariah, the son of Shephatiah, the son of Mahalaleel, of the children of Perez;

in Hierusalem habitauerunt de filiis Iuda et de filiis Beniamin de filiis Iuda Athaias filius Aziam filii Zacchariae filii

11:5 And Maaseiah the son of Baruch, the son of Col- Amariae filii Saphatia filii Malelehel de filiis Phares 11:5

hozeh, the son of Hazaiah, the son of Adaiah, the son of Joiarib, the son of Zechariah, the son of Shiloni.

Imaasia filius Baruch filius Coloza filius Azia filius Adaia

11:6 All the sons of Perez that dwelt at Jerusalem were filius Ioiarib filius Zacchariae filius Silonites 11:6 omnes fi-

four hundred threescore and eight valiant men.

lii Phares qui habitauerunt in Hierusalem quadringenti sexa-

11:7 And these are the sons of Benjamin; Sallu the son of ginta octo uiri fortes 11:7 hii sunt autem filii Beniamin Sel-

Meshullam, the son of Joed, the son of Pedaiah, the son of Kolaiah, the son of Maaseiah, the son of Ithiel, the son of

Jesaiah. lum filius Mosollam filius Ioed filius Phadaia filius Colaia 11:8 And after him Gabbai, Sallai, nine hundred twenty filius Masia filius Ethehel filius Isaia 11:8 et post eum Gab-

and eight.

11:9 And Joel the son of Zichri was their overseer: and bai Sellai nongenti uiginti octo 11:9 et Iohel filius Zechri

Judah the son of Senuah was second over the city.

praepositus eorum et Iuda filius Sennua super ciuitatem se-

11:10 Of the priests: Jedaiah the son of Joiarib, Jachin. cundus 11:10 et de sacerdotibus Idaia filius Ioarib Iachin 11:11 Seraiah the son of Hilkiah, the son of Meshullam, 11:11 Saraia filius Elcia filius Mesollam filius Sadoc filius

the son of Zadok, the son of Meraioth, the son of Ahitub, was the ruler of the house of God.

Meraioth filius Ahitub princeps domus Dei 11:12 et fratres

11:12 And their brethren that did the work of the house were eight hundred twenty and two: and Adaiah the son of Jeroham, the son of Pelaliah, the son of Amzi, the son of Zechariah, the son of Pashur, the son of Malchiah.

eorum facientes opera templi octingenti uiginti duo et Adaia filius Ieroam filius Felelia filius Amsi filius Zacchariae filius

11:13 And his brethren, chief of the fathers, two hundred Phessur filius Melchiae 11:13 et fratres eius principes pat-

forty and two: and Amashai the son of Azareel, the son of Ahasai, the son of Meshillemoth, the son of Immer,

rum ducenti quadraginta duo et Amassai filius Azrihel filius

11:14 And their brethren, mighty men of valour, an hun- Aazi filius Mosollamoth filius Emmer 11:14 et fratres eo-

dred twenty and eight: and their overseer was Zabdiel, the son of one of the great men.

rum potentes nimis centum uiginti octo et praepositus eorum

11:15 Also of the Levites: Shemaiah the son of Hashub, Zabdihel filius potentium 11:15 et de Leuitis Sebenia filius

the son of Azrikam, the son of Hashabiah, the son of Bunni;

Asob filius Azaricam filius Asabia filius Boni 11:16 et Sa-

11:16 And Shabbethai and Jozabad, of the chief of the Levites, had the oversight of the outward business of the house of God.

bathai et Iozabed super opera quae erant forinsecus in domo

Liber Nehemiae

Dei a principibus Leuitarum 11:17 et Mathania filius Micha 11:17 And Mattaniah the son of Micha, the son of Zabdi,

the son of Asaph, was the principal to begin the thanksgiving in prayer: and Bakbukiah the second among his brethren, and Abda the son of Shammua, the son of Galal, the son of Jeduthun.

filius Zebdaei filius Asaph princeps ad laudandum et confitendum in oratione et Becbecia secundus de fratribus eius et Abda filius Sammua filius Galal filius Idithun 11:18 omnes 11:18 All the Levites in the holy city were two hundred

fourscore and four.

Leuitae in ciuitate sancta ducenti octoginta quattuor 11:19 11:19 Moreover the porters, Akkub, Talmon, and their

brethren that kept the gates, were an hundred seventy and two.

et ianitores Accob Telmon et fratres eorum qui custodiebant ostia centum septuaginta duo 11:20 et reliqui ex Israhel sa- 11:20 And the residue of Israel, of the priests, and the

Levites, were in all the cities of Judah, every one in his inheritance.

cerdotes et Leuitae in uniuersis ciuitatibus Iuda unusquisque in possessione sua 11:21 et Nathinnei qui habitabant in Ofel 11:21 But the Nethinims dwelt in Ophel: and Ziha and

Gispa were over the Nethinims.

et Siaha et Gaspha de Nathinneis 11:22 et episcopus Leuita- 11:22 The overseer also of the Levites at Jerusalem was

Uzzi the son of Bani, the son of Hashabiah, the son of Mattaniah, the son of Micha. Of the sons of Asaph, the

rum in Hierusalem Azzi filius Bani filius Asabiae filius Mat- singers were over the business of the house of God. thaniae filius Michae de filiis Asaph cantores in ministerio domus Dei 11:23 praeceptum quippe regis super eos erat et 11:23 For it was the king’s commandment concerning

them, that a certain portion should be for the singers, due for every day.

ordo in cantoribus per dies singulos 11:24 et Fataia filius

11:24 And Pethahiah the son of Meshezabeel, of the children of Zerah the son of Judah, was at the king’s hand in all matters concerning the people.

Mesezebel de filiis Zera filii Iuda in manu regis iuxta omne uerbum populi 11:25 et in domibus per omnes regiones eo- 11:25 And for the villages, with their fields, some of the

children of Judah dwelt at Kirjatharba, and in the villages thereof, and at Dibon, and in the villages thereof, and at

rum de filiis Iuda habitauerunt in Cariatharbe et in filiabus Jekabzeel, and in the villages thereof, eius et in Dibon et in filiabus eius et in Capsel et in uiculis eius 11:26 et in Iesue et in Molada et in Bethfaleth 11:27 11:26 And at Jeshua, and at Moladah, and at Bethphelet,

11:27 And at Hazarshual, and at Beersheba, and in the villages thereof,

et in Asersual et in Bersabee et in filiabus eius 11:28 et in

11:28 And at Ziklag, and at Mekonah, and in the villages thereof,

Siceleg et in Mochona et in filiabus eius 11:29 et in Ain-

11:29 And at Enrimmon, and at Zareah, and at Jarmuth,

remmon et in Sara et in Irimuth 11:30 Zanoa Odollam et 11:30 Zanoah, Adullam, and in their villages, at Lachish, and the fields thereof, at Azekah, and in the villages thereof. And they dwelt from Beersheba unto the valley

uillis earum Lachis et regionibus eius Azeca et filiabus eius of Hinnom. et manserunt in Bersabee usque ad uallem Ennom 11:31 fi- 11:31 The children also of Benjamin from Geba dwelt at

Michmash, and Aija, and Bethel, and in their villages.

lii autem Beniamin a Geba Mechmas et Aia et Bethel et filiabus eius 11:32 Anathoth Nob Anania 11:33 Asor Rama 11:32 And at Anathoth, Nob, Ananiah,

11:33 Hazor, Ramah, Gittaim,

Getthaim 11:34 Adid Seboim Neballa Loth 11:35 et Ono _{11:34 Hadid, Zeboim, Neballat,}

11:35 Lod, and Ono, the valley of craftsmen.

ualle Artificum 11:36 et de Leuitis partitiones Iuda et Be- _{11:36 And of the Levites were divisions in Judah, and in}

Benjamin.

niamin

12:1 hii autem sacerdotes et Leuitae qui ascenderunt cum 12:1 Now these are the priests and the Levites that went up

with Zerubbabel the son of Shealtiel, and Jeshua: Seraiah, Jeremiah, Ezra,

Zorobabel filio Salathihel et Iosue Saraia Hieremias Ezra 12:2 Amaria Melluch Attus 12:3 Sechenia Reum Meremuth 12:2 Amariah, Malluch, Hattush,

12:3 Shechaniah, Rehum, Meremoth,

12:4 Addo Genthon Abia 12:5 Miamin Madia Belga 12:6 _{12:4 Iddo, Ginnetho, Abijah,}

12:5 Miamin, Maadiah, Bilgah, 12:6 Shemaiah, and Joiarib, Jedaiah,

Biblia Sacra Vulgata

12:7 Sallu, Amok, Hilkiah, Jedaiah. These were the chief Semaia et Ioarib Idaia Sellum Amoc Elceia 12:7 Idaia isti of the priests and of their brethren in the days of Jeshua.

12:8 Moreover the Levites: Jeshua, Binnui, Kadmiel, principes sacerdotum et fratres eorum in diebus Iosue 12:8

Sherebiah, Judah, and Mattaniah, which was over the thanksgiving, he and his brethren.

porro Leuitae Iesua Bennui Cedmihel Sarabia Iuda Mat-

12:9 Also Bakbukiah and Unni, their brethren, were over hanias super hymnos ipsi et fratres eorum 12:9 et Becbeagainst them in the watches.

cia atque et Hanni fratres eorum unusquisque in officio suo

12:10 And Jeshua begat Joiakim, Joiakim also begat 12:10 Hiesue autem genuit Ioachim et Ioachim genuit Elia-

Eliashib, and Eliashib begat Joiada,

12:11 And Joiada begat Jonathan, and Jonathan begat Jad- sib et Eliasib genuit Ioiada 12:11 et Ioiada genuit Ionat-

dua. 12:12 And in the days of Joiakim were priests, the chief of han et Ionathan genuit Ieddoa 12:12 in diebus autem Ioathe fathers: of Seraiah, Meraiah; of Jeremiah, Hananiah;

chim erant sacerdotes principes familiarum Saraiae Amaria

12:13 Of Ezra, Meshullam; of Amariah, Jehohanan; Hieremiae Anania 12:13 Ezrae Mosollam Amariae Iohanan 12:14 Of Melicu, Jonathan; of Shebaniah, Joseph; 12:14 Milico Ionathan Sebeniae Ioseph 12:15 Arem Edna

12:15 Of Harim, Adna; of Meraioth, Helkai;

12:16 Of Iddo, Zechariah; of Ginnethon, Meshullam; Maraioth Elci 12:16 Addaiae Zaccharia Genthon Mosollam

12:17 Of Abijah, Zichri; of Miniamin, of Moadiah, Piltai: 12:17 Abiae Zecheri Miamin et Moadiae Felti 12:18 Belgae 12:18 Of Bilgah, Shammua; of Shemaiah, Jehonathan;

12:19 And of Joiarib, Mattenai; of Jedaiah, Uzzi; Sammua Semaiae Ionathan 12:19 Ioiarib Matthanai Iadaiae 12:20 Of Sallai, Kallai; of Amok, Eber; Azzi 12:20 Sellaiae Celai Amoc Eber 12:21 Elciae Asebia

12:21 Of Hilkiah, Hashabiah; of Jedaiah, Nethaneel.

12:22 The Levites in the days of Eliashib, Joiada, and Jo- Idaiae Nathanahel 12:22 Leuitae in diebus Eliasib et Ioiada

hanan, and Jaddua, were recorded chief of the fathers: also the priests, to the reign of Darius the Persian.

et Ionan et Ieddoa scripti principes familiarum et sacerdotes

12:23 The sons of Levi, the chief of the fathers, were writ- in regno Darii Persae 12:23 filii Leui principes familiarum

ten in the book of the chronicles, even until the days of Johanan the son of Eliashib.

scripti in libro uerborum dierum et usque ad dies Ionathan

12:24 And the chief of the Levites: Hashabiah, Sherebiah, filii Eliasib 12:24 et principes Leuitarum Asebia Serebia et

and Jeshua the son of Kadmiel, with their brethren over against them, to praise and to give thanks, according to the commandment of David the man of God, ward over against ward.

Iesue filius Cedmihel et fratres eorum per uices suas ut laudarent et confiterentur iuxta praeceptum Dauid uiri Dei et

12:25 Mattaniah, and Bakbukiah, Obadiah, Meshullam, obseruarent aeque per ordinem 12:25 Matthania et Becbe-

Talmon, Akkub, were porters keeping the ward at the thresholds of the gates.

cia Obedia Mosollam Thelmon Accub custodes portarum et

12:26 These were in the days of Joiakim the son of Jeshua, uestibulorum ante portas 12:26 hii in diebus Ioachim filii

the son of Jozadak, and in the days of Nehemiah the governor, and of Ezra the priest, the scribe.

Iesue filii Iosedech et in diebus Neemiae ducis et Ezrae sa-

12:27 And at the dedication of the wall of Jerusalem they cerdotis scribaeque 12:27 in dedicatione autem muri Hie-

sought the Levites out of all their places, to bring them to Jerusalem, to keep the dedication with gladness, both with thanksgivings, and with singing, with cymbals, psalteries, and with harps.

rusalem requisierunt Leuitas de omnibus locis suis ut adducerent eos in Hierusalem et facerent dedicationem et laetitiam in actione gratiarum et in cantico in cymbalis psalteriis

12:28 And the sons of the singers gathered themselves et citharis 12:28 congregati sunt ergo filii cantorum et de

together, both out of the plain country round about Jerusalem, and from the villages of Netophathi;

campestribus circa Hierusalem et de uillis Netuphati 12:29

12:29 Also from the house of Gilgal, and out of the fields of Geba and Azmaveth: for the singers had builded them villages round about Jerusalem.

et de domo Galgal et de regionibus Geba et Azmaueth quo-

Liber Nehemiae

niam uillas aedificauerunt sibi cantores in circuitu Hierusa-

lem 12:30 et mundati sunt sacerdotes et Leuitae et munda- 12:30 And the priests and the Levites purified themselves,

and purified the people, and the gates, and the wall.

uerunt populum et portas et murum 12:31 ascendere autem 12:31 Then I brought up the princes of Judah upon the

wall, and appointed two great companies of them that gave thanks, whereof one went on the right hand upon the wall

feci principes Iuda super murum et statui duos choros lau- toward the dung gate: dantium magnos et ierunt ad dexteram super murum ad portam Sterquilinii 12:32 et iuit post eos Osaias et media pars 12:32 And after them went Hoshaiah, and half of the

princes of Judah,

principum Iuda 12:33 et Azarias Ezras et Mosollam Iuda 12:33 And Azariah, Ezra, and Meshullam, Judah, and

Benjamin, and Shemaiah, and Jeremiah,

et Beniamin et Semeia et Hieremia 12:34 et de filiis sacer- 12:34 And certain of the priests’ sons with trumpets;

namely, Zechariah the son of Jonathan, the son of Shemaiah, the son of Mattaniah, the son of Michaiah, the son

dotum in tubis Zaccharias filius Ionathan filius Semeiae fi- of Zaccur, the son of Asaph: lius Mathaniae filius Michaiae filius Zecchur filius Asaph 12:35 et fratres eius Semeia et Azarel Malalai Galalai Maai 12:35 And his brethren, Shemaiah, and Azarael, Milalai,

Gilalai, Maai, Nethaneel, and Judah, Hanani, with the musical instruments of David the man of God, and Ezra the

Nathanel et Iuda et Anani in uasis cantici Dauid uiri Dei scribe before them. et Ezras scriba ante eos in porta Fontis 12:36 et contra eos 12:36 And at the fountain gate, which was over against

them, they went up by the stairs of the city of David, at the going up of the wall, above the house of David, even unto

ascenderunt in gradibus ciuitatis Dauid in ascensu muri su- the water gate eastward. per domum Dauid et usque ad portam Aquarum ad orientem 12:37 et chorus secundus gratias referentium ibat ex aduerso 12:37 And the other company of them that gave thanks

went over against them, and I after them, and the half of the people upon the wall, from beyond the tower of the

et ego post eum et media pars populi super murum et super furnaces even unto the broad wall; turrem Furnorum et usque ad murum latissimum 12:38 et 12:38 And from above the gate of Ephraim, and above the

old gate, and above the fish gate, and the tower of Hananeel, and the tower of Meah, even unto the sheep gate: and

super portam Ephraim et super portam Antiquam et super they stood still in the prison gate. portam Piscium et turrem Ananehel et turrem Ema et usque ad portam Gregis et steterunt in porta Custodiae 12:39 ste- 12:39 So stood the two companies of them that gave

thanks in the house of God, and I, and the half of the rulers with me:

teruntque duo chori laudantium in domo Dei et ego et dimidia pars magistratuum mecum 12:40 et sacerdotes Eliachim 12:40 And the priests; Eliakim, Maaseiah, Miniamin,

Michaiah, Elioenai, Zechariah, and Hananiah, with trumpets;

Maasia Miniamin Michea Elioenai Zaccharia Anania in tubis 12:41 et Maasia et Semea et Eleazar et Azi et Iohanan 12:41 And Maaseiah, and Shemaiah, and Eleazar, and

Uzzi, and Jehohanan, and Malchijah, and Elam, and Ezer. And the singers sang loud, with Jezrahiah their overseer.

et Melchia et Elam et Ezer et clare cecinerunt cantores et Iezraia praepositus 12:42 et immolauerunt in die illa uicti- 12:42 Also that day they offered great sacrifices, and re-

joiced: for God had made them rejoice with great joy: the wives also and the children rejoiced: so that the joy of

mas magnas et laetati sunt Deus enim laetificauerat eos lae- Jerusalem was heard even afar off. titia magna sed et uxores eorum et liberi gauisi sunt et audita est laetitia Hierusalem procul 12:43 recensuerunt quoque in 12:43 And at that time were some appointed over the

chambers for the treasures, for the offerings, for the firstfruits, and for the tithes, to gather into them out of the fields of the cities the portions of the law for the priests and Levites: for Judah rejoiced for the priests and for the Levites that waited.

die illa uiros super gazofilacia thesauri ad libamina et ad primitias et ad decimas ut introferrent per eos principes ciuitatis in decore gratiarum actionis sacerdotes et Leuitas quia lae-

Biblia Sacra Vulgata

12:44 And both the singers and the porters kept the ward tatus est Iuda in sacerdotibus et Leuitis adstantibus 12:44

of their God, and the ward of the purification, according to the commandment of David, and of Solomon his son.

et custodierunt obseruationem Dei sui et obseruationem expiationis et cantores et ianitores iuxta praeceptum Dauid et

12:45 For in the days of David and Asaph of old there were Salomonis filii eius 12:45 quia in diebus Dauid et Asaph ab

chief of the singers, and songs of praise and thanksgiving unto God.

exordio erant principes constituti cantorum in carmine lau-

12:46 And all Israel in the days of Zerubbabel, and in the dantium et confitentium Deo 12:46 et omnis Israhel in die-

days of Nehemiah, gave the portions of the singers and the porters, every day his portion: and they sanctified holy things unto the Levites; and the Levites sanctified them unto the children of Aaron.

bus Zorobabel et in diebus Neemiae dabat partes cantoribus et ianitoribus per dies singulos et sanctificabant Leuitas et Leuitae sanctificabant filios Aaron

13:1 On that day they read in the book of Moses in the 13:1 in die autem illo lectum est in uolumine Mosi au-

audience of the people; and therein was found written, that the Ammonite and the Moabite should not come into the

congregation of God for ever; diente populo et inuentum est scriptum in eo quod non de-

beat introire Ammanites et Moabites in ecclesiam Dei us-

13:2 Because they met not the children of Israel with bread que in aeternum 13:2 eo quod non occurrerint filiis Israhel and with water, but hired Balaam against them, that he should curse them: howbeit our God turned the curse into

a blessing. cum pane et aqua et conduxerint aduersum eum Balaam ad

maledicendum ei et conuertit Deus noster maledictionem in

13:3 Now it came to pass, when they had heard the law, benedictionem 13:3 factum est autem cum audissent legem

that they separated from Israel all the mixed multitude. 13:4 And before this, Eliashib the priest, having the over- separauerunt omnem alienigenam ab Israhel 13:4 et super sight of the chamber of the house of our God, was allied unto Tobiah:

hoc erat Eliasib sacerdos qui fuerat positus in gazofilacio

13:5 And he had prepared for him a great chamber, where domus Dei nostri et proximus Tobiae 13:5 fecit ergo sibi

aforetime they laid the meat offerings, the frankincense, and the vessels, and the tithes of the corn, the new wine, and the oil, which was commanded to be given to the Levites, and the singers, and the porters; and the offerings of the priests.

gazofilacium grande et ibi erant ante eum reponentes munera et tus et uasa et decimam frumenti et uini et olei partes Leuitarum et cantorum et ianitorum et primitias sacer-

13:6 But in all this time was not I at Jerusalem: for in the dotales 13:6 in omnibus autem his non fui in Hierusalem two and thirtieth year of Artaxerxes king of Babylon came I unto the king, and after certain days obtained I leave of

the king: quia in anno tricesimo secundo Artarxersis regis Babylonis

13:7 And I came to Jerusalem, and understood of the evil ueni ad regem et in fine dierum rogaui regem 13:7 et ueni that Eliashib did for Tobiah, in preparing him a chamber in the courts of the house of God.

in Hierusalem et intellexi malum quod fecerat Eliasib To-

13:8 And it grieved me sore: therefore I cast forth all the biae ut faceret ei thesaurum in uestibulis domus Dei 13:8 household stuff to Tobiah out of the chamber.

et malum mihi uisum est ualde et proieci uasa domus To-

13:9 Then I commanded, and they cleansed the chambers: biae foras de gazofilacio 13:9 praecepique et mundauerunt

and thither brought I again the vessels of the house of God, with the meat offering and the frankincense.

gazofilacia et rettuli ibi uasa domus Dei sacrificium et tus

13:10 And I perceived that the portions of the Levites had 13:10 et cognoui quoniam partes Leuitarum non fuissent da-

not been given them: for the Levites and the singers, that did the work, were fled every one to his field.

tae et fugisset unusquisque in regionem suam de Leuitis et

13:11 Then contended I with the rulers, and said, Why is de cantoribus et de his qui ministrabant 13:11 et egi cau-

the house of God forsaken? And I gathered them together, and set them in their place.

Liber Nehemiae

sam aduersus magistratus et dixi quare dereliquimus domum Dei et congregaui eos et feci stare in stationibus suis 13:12 13:12 Then brought all Judah the tithe of the corn and the

new wine and the oil unto the treasuries.

et omnis Iuda adportabat decimam frumenti et uini et olei in horrea 13:13 et constituimus super horrea Selemiam sa- 13:13 And I made treasurers over the treasuries,

Shelemiah the priest, and Zadok the scribe, and of the Levites, Pedaiah: and next to them was Hanan the son of Zaccur, the son of Mattaniah: for they were counted faithful, and their office was to distribute unto their brethren.

cerdotem et Sadoc scribam et Phadaiam de Leuitis et iuxta eos Anan filium Zacchur filium Matthaniae quoniam fideles conprobati sunt et ipsis creditae sunt partes fratrum suorum 13:14 memento mei Deus meus pro hoc et ne deleas mise- 13:14 Remember me, O my God, concerning this, and

wipe not out my good deeds that I have done for the house of my God, and for the offices thereof.

rationes meas quas feci in domo Dei mei et in caerimoniis eius 13:15 in diebus illis uidi in Iuda calcabant torcularia in 13:15 In those days saw I in Judah some treading wine

presses on the sabbath, and bringing in sheaves, and lading asses; as also wine, grapes, and figs, and all manner of burdens, which they brought into Jerusalem on the sabbath day: and I testified against them in the day wherein they sold victuals.

sabbato portantes aceruos et onerantes super asinos uinum et uuas et ficus et omne onus et inferentes Hierusalem in die sabbati et contestatus sum ut in die qua uendere liceret uenderent 13:16 et Tyrii habitauerunt in ea inferentes pisces et 13:16 There dwelt men of Tyre also therein, which brought

fish, and all manner of ware, and sold on the sabbath unto the children of Judah, and in Jerusalem.

omnia uenalia et uendebant in sabbatis filiis Iuda et in Hierusalem 13:17 et obiurgaui optimates Iuda et dixi eis quae 13:17 Then I contended with the nobles of Judah, and said

unto them, What evil thing is this that ye do, and profane the sabbath day?

est res haec mala quam uos facitis et profanatis diem sabbati 13:18 numquid non haec fecerunt patres nostri et adduxit De- 13:18 Did not your fathers thus, and did not our God bring

all this evil upon us, and upon this city? yet ye bring more wrath upon Israel by profaning the sabbath.

us noster super nos omne malum hoc et super ciuitatem hanc et uos additis iracundiam super Israhel uiolando sabbatum 13:19 factum est itaque cum quieuissent portae Hierusalem 13:19 And it came to pass, that when the gates of

Jerusalem began to be dark before the sabbath, I commanded that the gates should be shut, and charged that they should not be opened till after the sabbath: and some of my servants set I at the gates, that there should no burden be brought in on the sabbath day.

die sabbati dixi et cluserunt ianuas et praecepi ut non aperirent eas usque post sabbatum et de pueris meis constitui super portas ut nullus inferret onus in die sabbati 13:20 et 13:20 So the merchants and sellers of all kind of ware

lodged without Jerusalem once or twice.

manserunt negotiatores et uendentes uniuersa uenalia foris Hierusalem semel et bis 13:21 et contestatus sum eos et dixi 13:21 Then I testified against them, and said unto them,

Why lodge ye about the wall? if ye do so again, I will lay hands on you. From that time forth came they no more on

eis quare manetis ex aduerso muri si secundo hoc feceritis the sabbath. manum mittam in uos itaque ex tempore illo non uenerunt in sabbato 13:22 dixi quoque Leuitis ut mundarentur et ue- 13:22 And I commanded the Levites that they should

cleanse themselves, and that they should come and keep the gates, to sanctify the sabbath day. Remember me, O my God, concerning this also, and spare me according to the greatness of thy mercy.

nirent ad custodiendas portas et sanctificandum diem sabbati et pro hoc ergo memento mei Deus meus et parce mihi secundum multitudinem miserationum tuarum 13:23 sed et in 13:23 In those days also saw I Jews that had married wives

of Ashdod, of Ammon, and of Moab:

diebus illis uidi Iudaeos ducentes uxores azotias ammaniti-

Biblia Sacra Vulgata

13:24 And their children spake half in the speech of Ash- das et moabitidas 13:24 et filii eorum ex media parte lodod, and could not speak in the Jews’ language, but according to the language of each people.

quebantur azotice et nesciebant loqui iudaice et loqueban-

13:25 And I contended with them, and cursed them, and tur iuxta linguam populi et populi 13:25 et obiurgaui eos

smote certain of them, and plucked off their hair, and made them swear by God, saying, Ye shall not give your daughters unto their sons, nor take their daughters unto your sons, or for yourselves.

et maledixi et cecidi ex ipsis uiros et decaluaui eos et adiuraui in Deo ut non darent filias suas filiis eorum et non acciperent de filiabus eorum filiis suis et sibimet ipsis dicens

13:26 Did not Solomon king of Israel sin by these things? 13:26 numquid non in huiuscemodi re peccauit Salomon rex

yet among many nations was there no king like him, who was beloved of his God, and God made him king over all Israel: nevertheless even him did outlandish women cause to sin.

Israhel et certe in gentibus multis non erat rex similis ei et dilectus Deo suo erat et posuit eum Deus regem super omnem Israhel et ipsum ergo ad peccatum duxerunt mulieres

13:27 Shall we then hearken unto you to do all this great alienigenae 13:27 numquid et nos inoboedientes faciemus

evil, to transgress against our God in marrying strange wives?

omne malum grande hoc ut praeuaricemur in Deo nostro et

13:28 And one of the sons of Joiada, the son of Eliashib ducamus uxores peregrinas 13:28 de filiis autem Ioiada fi-

the high priest, was son in law to Sanballat the Horonite: therefore I chased him from me.

lii Eliasib sacerdotis magni gener erat Sanaballat Horonitis

13:29 Remember them, O my God, because they have de- quem fugaui a me 13:29 recordare Domine Deus meus ad-

filed the priesthood, and the covenant of the priesthood, and of the Levites.

uersum eos qui polluunt sacerdotium iusque sacerdotale et

13:30 Thus cleansed I them from all strangers, and ap- leuiticum 13:30 igitur mundaui eos ab omnibus alienigenis

pointed the wards of the priests and the Levites, every one in his business;

et constitui ordines sacerdotum et Leuitarum unumquemque

13:31 And for the wood offering, at times appointed, and in ministerio suo 13:31 et in oblatione lignorum in tempo-

for the firstfruits. Remember me, O my God, for good.

\underline{Liber Esther}

ribus constitutis et in primitiis memento mei Deus meus in bonum 1:1 in diebus Asueri qui regnauit ab India usque Aethio- 1:1 Now it came to pass in the days of Ahasuerus, (this is

Ahasuerus which reigned, from India even unto Ethiopia, over an hundred and seven and twenty provinces:)

piam super centum uiginti septem prouincias 1:2 quando

1:2 That in those days, when the king Ahasuerus sat on the throne of his kingdom, which was in Shushan the palace,

sedit in solio regni sui Susa ciuitas regni eius exordium fuit 1:3 tertio igitur anno imperii sui fecit grande conui- 1:3 In the third year of his reign, he made a feast unto all

his princes and his servants; the power of Persia and Media, the nobles and princes of the provinces, being before

uium cunctis principibus et pueris suis fortissimis Persa- him: rum et Medorum inclitis et praefectis prouinciarum coram se 1:4 ut ostenderet diuitias gloriae regni sui ac magnitudi- 1:4 When he shewed the riches of his glorious kingdom

and the honour of his excellent majesty many days, even an hundred and fourscore days.

nem atque iactantiam potentiae suae multo tempore centum uidelicet et octoginta diebus 1:5 cumque implerentur dies 1:5 And when these days were expired, the king made a

feast unto all the people that were present in Shushan the palace, both unto great and small, seven days, in the court

conuiuii inuitauit omnem populum qui inuentus est Susis of the garden of the king’s palace; a maximo usque ad minimum et septem diebus iussit conuiuium praeparari in uestibulo horti et nemoris quod regio

cultu et manu consitum erat 1:6 et pendebant ex omni parte 1:6 Where were white, green, and blue, hangings, fastened with cords of fine linen and purple to silver rings and pillars of marble: the beds were of gold and silver, upon a

tentoria aerii coloris et carpasini et hyacinthini sustentata fu- pavement of red, and blue, and white, and black, marble. nibus byssinis atque purpureis qui eburneis circulis inserti erant et columnis marmoreis fulciebantur lectuli quoque aurei et argentei super pauimentum zmaragdino et pario stratum lapide dispositi erant quod mira uarietate pictura decorabat 1:7 bibebant autem qui inuitati erant aureis poculis et 1:7 And they gave them drink in vessels of gold, (the ves-

sels being diverse one from another,) and royal wine in abundance, according to the state of the king.

aliis atque aliis uasis cibi inferebantur uinum quoque ut magnificentia regia dignum erat abundans et praecipuum ponebatur 1:8 nec erat qui nolentes cogeret ad bibendum sed sic 1:8 And the drinking was according to the law; none did

compel: for so the king had appointed to all the officers of his house, that they should do according to every man’s

rex statuerat praeponens mensis singulos de principibus suis pleasure. ut sumeret unusquisque quod uellet 1:9 Vasthi quoque re- 1:9 Also Vashti the queen made a feast for the women in

the royal house which belonged to king Ahasuerus.

gina fecit conuiuium feminarum in palatio ubi rex Asuerus manere consueuerat 1:10 itaque die septimo cum rex esset 1:10 On the seventh day, when the heart of the king was

merry with wine, he commanded Mehuman, Biztha, Harbona, Bigtha, and Abagtha, Zethar, and Carcas, the seven chamberlains that served in the presence of Ahasuerus the king,

hilarior et post nimiam potionem incaluisset mero praecepit Mauman et Bazatha et Arbona et Bagatha et Abgatha et Zarath et Charchas septem eunuchis qui in conspectu eius ministrabant 1:11 ut introducerent reginam Vasthi coram re- 1:11 To bring Vashti the queen before the king with the

crown royal, to shew the people and the princes her beauty: for she was fair to look on.

ge posito super caput eius diademate et ostenderet cunctis populis et principibus illius pulchritudinem erat enim pulchra ualde 1:12 quae rennuit et ad regis imperium quod 1:12 But the queen Vashti refused to come at the king’s

commandment by his chamberlains: therefore was the king very wroth, and his anger burned in him.

Biblia Sacra Vulgata

per eunuchos mandauerat uenire contempsit unde iratus rex

1:13 Then the king said to the wise men, which knew the et nimio furore succensus 1:13 interrogauit sapientes qui

times, (for so was the king’s manner toward all that knew law and judgment:

ex more regio semper ei aderant et illorum faciebat cuncta

1:14 And the next unto him was Carshena, Shethar, Ad- consilio scientium leges ac iura maiorum 1:14 erant autem

matha, Tarshish, Meres, Marsena, and Memucan, the seven princes of Persia and Media, which saw the king’s

face, and which sat the first in the kingdom;) primi et proximi Charsena et Sethar et Admatha et Tharsis et Mares et Marsana et Mamucha septem duces Persarum atque Medorum qui uidebant faciem regis et primi post eum

1:15 What shall we do unto the queen Vashti according to residere soliti erant 1:15 cui sententiae Vasthi regina subia-

law, because she hath not performed the commandment of the king Ahasuerus by the chamberlains?

ceret quae Asueri regis imperium quod per eunuchos man-

1:16 And Memucan answered before the king and the dauerat facere noluisset 1:16 responditque Mamuchan au-

princes, Vashti the queen hath not done wrong to the king only, but also to all the princes, and to all the people that

are in all the provinces of the king Ahasuerus. diente rege atque principibus non solum regem laesit regina

Vasthi sed omnes principes et populos qui sunt in cunctis

1:17 For this deed of the queen shall come abroad unto all prouinciis regis Asueri 1:17 egredietur enim sermo reginae

women, so that they shall despise their husbands in their eyes, when it shall be reported, The king Ahasuerus commanded Vashti the queen to be brought in before him, but she came not.

ad omnes mulieres ut contemnant uiros suos et dicant rex Asuerus iussit ut regina Vasthi intraret ad eum et illa noluit

1:18 Likewise shall the ladies of Persia and Media say this 1:18 atque hoc exemplo omnes principum coniuges Persarum day unto all the king’s princes, which have heard of the deed of the queen. Thus shall there arise too much con-

tempt and wrath. atque Medorum paruipendent imperia maritorum unde regis 1:19 If it please the king, let there go a royal command- iusta est indignatio 1:19 et si tibi placet egrediatur edictum

ment from him, and let it be written among the laws of the Persians and the Medes, that it be not altered, That Vashti come no more before king Ahasuerus; and let the king give her royal estate unto another that is better than she.

a facie tua et scribatur iuxta legem Persarum atque Medorum quam praeteriri inlicitum est ut nequaquam ultra Vasthi ingrediatur ad regem sed regnum illius altera quae melior

1:20 And when the king’s decree which he shall make illa est accipiat 1:20 et hoc in omne quod latissimum est

shall be published throughout all his empire, (for it is great,) all the wives shall give to their husbands honour,

both to great and small. prouinciarum tuarum diuulgetur imperium et cunctae uxo-

1:21 And the saying pleased the king and the princes; and res tam maiorum quam minorum deferant maritis suis 1:21 the king did according to the word of Memucan:

placuit consilium eius regi et principibus fecitque rex iuxta

1:22 For he sent letters into all the king’s provinces, into consultum Mamuchan 1:22 et misit epistulas ad uniuersas every province according to the writing thereof, and to every people after their language, that every man should bear rule in his own house, and that it should be published according to the language of every people.

prouincias regni sui ut quaeque gens audire et legere poterat diuersis linguis et litteris esse uiros principes ac maiores in domibus suis et hoc per cunctos populos diuulgari

2:1 After these things, when the wrath of king Ahasuerus 2:1 his itaque gestis postquam regis Asueri deferbuerat in-

was appeased, he remembered Vashti, and what she had done, and what was decreed against her.

dignatio recordatus est Vasthi et quae fecisset uel quae passa

2:2 Then said the king’s servants that ministered unto him, esset 2:2 dixeruntque pueri regis ac ministri eius quaeran-

Let there be fair young virgins sought for the king:

2:3 And let the king appoint officers in all the provinces tur regi puellae uirgines ac speciosae 2:3 et mittantur qui

of his kingdom, that they may gather together all the fair young virgins unto Shushan the palace, to the house of the women, unto the custody of Hege the king’s chamberlain, keeper of the women; and let their things for purification be given them:

DCCCXLII

considerent per uniuersas prouincias puellas speciosas et uirgines et adducant eas ad ciuitatem Susan et tradant in domum feminarum sub manu Aegaei eunuchi qui est praepositus et custos mulierum regiarum et accipiant mundum muliebrem et cetera ad usus necessaria 2:4 et quaecumque inter 2:4 And let the maiden which pleaseth the king be queen

instead of Vashti. And the thing pleased the king; and he did so.

omnes oculis regis placuerit ipsa regnet pro Vasthi placuit sermo regi et ita ut suggesserant iussit fieri 2:5 erat uir iu- 2:5 Now in Shushan the palace there was a certain Jew,

whose name was Mordecai, the son of Jair, the son of Shimei, the son of Kish, a Benjamite;

daeus in Susis ciuitate uocabulo Mardocheus filius Iair filii Semei filii Cis de stirpe Iemini 2:6 qui translatus fuerat de 2:6 Who had been carried away from Jerusalem with the

captivity which had been carried away with Jeconiah king of Judah, whom Nebuchadnezzar the king of Babylon had

Hierusalem eo tempore quo Iechoniam regem Iuda Nabu- carried away. chodonosor rex Babylonis transtulerat 2:7 qui fuit nutricius 2:7 And he brought up Hadassah, that is, Esther, his un-

cle’s daughter: for she had neither father nor mother, and the maid was fair and beautiful; whom Mordecai, when her father and mother were dead, took for his own daughter.

filiae fratris sui Edessae quae altero nomine Hester uocabatur et utrumque parentem amiserat pulchra nimis et decora facie mortuisque patre eius ac matre Mardocheus sibi eam adoptauit in filiam 2:8 cumque percrebuisset regis impe- 2:8 So it came to pass, when the king’s commandment and

his decree was heard, and when many maidens were gathered together unto Shushan the palace, to the custody of Hegai, that Esther was brought also unto the king’s house, to the custody of Hegai, keeper of the women.

rium et iuxta mandata illius multae uirgines pulchrae adducerentur Susan et Aegaeo traderentur eunucho Hester quoque inter ceteras puellas ei tradita est ut seruaretur in numero feminarum 2:9 quae placuit ei et inuenit gratiam in 2:9 And the maiden pleased him, and she obtained kind-

ness of him; and he speedily gave her her things for purification, with such things as belonged to her, and seven maidens, which were meet to be given her, out of the king’s house: and he preferred her and her maids unto the best place of the house of the women.

conspectu illius ut adceleraret mundum muliebrem et traderet ei partes suas et septem puellas speciosissimas de domo regis et tam ipsam quam pedisequas eius ornaret atque excoleret 2:10 quae noluit indicare ei populum et patriam suam 2:10 Esther had not shewed her people nor her kindred:

for Mordecai had charged her that she should not shew it.

Mardocheus enim praeceperat ut de hac re omnino reticeret 2:11 qui deambulabat cotidie ante uestibulum domus in 2:11 And Mordecai walked every day before the court of

the women’s house, to know how Esther did, and what should become of her.

qua electae uirgines seruabantur curam agens salutis Hester et scire uolens quid ei accideret 2:12 cum autem uenis- 2:12 Now when every maid’s turn was come to go in to

king Ahasuerus, after that she had been twelve months, according to the manner of the women, (for so were the days of their purifications accomplished, to wit, six months with oil of myrrh, and six months with sweet odours, and with other things for the purifying of the women;)

set tempus singularum per ordinem puellarum ut intrarent ad regem expletis omnibus quae ad cultum muliebrem pertinebant mensis duodecimus uertebatur ita dumtaxat ut sex menses oleo unguerentur myrtino et aliis sex quibusdam pigmentis et aromatibus uterentur 2:13 ingredientesque ad re- 2:13 Then thus came every maiden unto the king; what-

soever she desired was given her to go with her out of the house of the women unto the king’s house.

gem quicquid postulassent ad ornatum pertinens accipiebant

Biblia Sacra Vulgata

et ut eis placuerat conpositae de triclinio feminarum ad regis

2:14 In the evening she went, and on the morrow she re- cubiculum transiebant 2:14 et quae intrauerat uespere eg-

turned into the second house of the women, to the custody of Shaashgaz, the king’s chamberlain, which kept the concubines: she came in unto the king no more, except the king delighted in her, and that she were called by name.

rediebatur mane atque inde in secundas aedes deducebatur quae sub manu Sasagazi eunuchi erant qui concubinis regis praesidebat nec habebat potestatem ad regem ultra redeendi

2:15 Now when the turn of Esther, the daughter of Abihail nisi uoluisset rex et eam uenire iussisset ex nomine 2:15

the uncle of Mordecai, who had taken her for his daughter, was come to go in unto the king, she required nothing but what Hegai the king’s chamberlain, the keeper of the women, appointed. And Esther obtained favour in the sight of all them that looked upon her.

euoluto autem tempore per ordinem instabat dies quo Hester filia Abiahil fratris Mardochei quam sibi adoptauerat in filiam intrare deberet ad regem quae non quaesiuit muliebrem cultum sed quaecumque uoluit Aegaeus eunuchus custos uirginum haec ei ad ornatum dedit erat enim formonsa ualde et incredibili pulchritudine omnium oculis gratiosa et

2:16 So Esther was taken unto king Ahasuerus into his amabilis uidebatur 2:16 ducta est itaque ad cubiculum re-

house royal in the tenth month, which is the month Tebeth, in the seventh year of his reign.

gis Asueri mense decimo qui uocatur tebeth septimo anno

2:17 And the king loved Esther above all the women, and regni eius 2:17 et amauit eam rex plus quam omnes mulie-

she obtained grace and favour in his sight more than all the virgins; so that he set the royal crown upon her head,

and made her queen instead of Vashti. res habuitque gratiam et misericordiam coram eo super om-

nes mulieres et posuit diadema regni in capite eius fecitque

2:18 Then the king made a great feast unto all his princes eam regnare in loco Vasthi 2:18 et iussit conuiuium prae-

and his servants, even Esther’s feast; and he made a release to the provinces, and gave gifts, according to the state of

the king. parari permagnificum cunctis principibus et seruis suis pro coniunctione et nuptiis Hester et dedit requiem in uniuersis prouinciis ac dona largitus est iuxta magnificentiam prin-

2:19 And when the virgins were gathered together the sec- cipalem 2:19 cumque et secundo quaererentur uirgines et ond time, then Mordecai sat in the king’s gate.

2:20 Esther had not yet shewed her kindred nor her peo- congregarentur Mardocheus manebat ad regis ianuam 2:20

ple; as Mordecai had charged her: for Esther did the commandment of Mordecai, like as when she was brought up

with him. necdumque prodiierat Hester patriam et populum suum iuxta mandatum eius quicquid enim ille praecipiebat obseruabat Hester et ita cuncta faciebat ut eo tempore solita erat quo

2:21 In those days, while Mordecai sat in the king’s gate, eam paruulam nutriebat 2:21 eo igitur tempore quo Martwo of the king’s chamberlains, Bigthan and Teresh, of those which kept the door, were wroth, and sought to lay

hands on the king Ahasuerus. docheus ad regis ianuam morabatur irati sunt Bagathan et Thares duo eunuchi regis qui ianitores erant et in primo palatii limine praesidebant uolueruntque insurgere in regem et

2:22 And the thing was known to Mordecai, who told it occidere eum 2:22 quod Mardocheum non latuit statimque

unto Esther the queen; and Esther certified the king thereof in Mordecai’s name.

nuntiauit reginae Hester et illa regi ex nomine Mardochei

2:23 And when inquisition was made of the matter, it was qui ad se rem detulerat 2:23 quaesitum est et inuentum et

found out; therefore they were both hanged on a tree: and it was written in the book of the chronicles before the king.

adpensus uterque eorum in patibulo mandatumque historiis et annalibus traditum coram rege

3:1 post haec rex Asuerus exaltauit Aman filium Amadathi 3:1 After these things did king Ahasuerus promote Haman

the son of Hammedatha the Agagite, and advanced him, and set his seat above all the princes that were with him.

qui erat de stirpe Agag et posuit solium eius super omnes

principes quos habebat 3:2 cunctique serui regis qui in fo- 3:2 And all the king’s servants, that were in the king’s gate, bowed, and reverenced Haman: for the king had so commanded concerning him. But Mordecai bowed not, nor

ribus palatii uersabantur flectebant genu et adorabant Aman did him reverence. sic enim eis praeceperat imperator solus Mardocheus non flectebat genu neque adorabat eum 3:3 cui dixerunt regis 3:3 Then the king’s servants, which were in the king’s

gate, said unto Mordecai, Why transgressest thou the king’s commandment?

pueri qui ad fores palatii praesidebant cur praeter ceteros non obseruas mandata regis 3:4 cumque hoc crebrius dicerent 3:4 Now it came to pass, when they spake daily unto him,

and he hearkened not unto them, that they told Haman, to see whether Mordecai’s matters would stand: for he had

et ille nollet audire nuntiauerunt Aman scire cupientes ut- told them that he was a Jew. rum perseueraret in sententia dixerat enim eis se esse Iudaeum 3:5 quod cum audisset Aman et experimento pro- 3:5 And when Haman saw that Mordecai bowed not, nor

did him reverence, then was Haman full of wrath.

basset quod Mardocheus non sibi flecteret genu nec se adoraret iratus est ualde 3:6 et pro nihilo duxit in unum Mar- 3:6 And he thought scorn to lay hands on Mordecai alone;

for they had shewed him the people of Mordecai: wherefore Haman sought to destroy all the Jews that were throughout the whole kingdom of Ahasuerus, even the people of Mordecai.

docheum mittere manus suas audierat enim quod esset gentis iudaeae magisque uoluit omnem Iudaeorum qui erant in regno Asueri perdere nationem 3:7 mense primo cuius uoca- 3:7 In the first month, that is, the month Nisan, in the

twelfth year of king Ahasuerus, they cast Pur, that is, the lot, before Haman from day to day, and from month to

bulum est nisan anno duodecimo regni Asueri missa est sors month, to the twelfth month, that is, the month Adar. in urnam quae hebraice dicitur phur coram Aman quo die et quo mense gens Iudaeorum deberet interfici et exiuit mensis duodecimus qui uocatur adar 3:8 dixitque Aman regi 3:8 And Haman said unto king Ahasuerus, There is a cer-

tain people scattered abroad and dispersed among the people in all the provinces of thy kingdom; and their laws are diverse from all people; neither keep they the king’s laws: therefore it is not for the king’s profit to suffer them.

Asuero est populus per omnes prouincias regni tui dispersus et a se mutuo separatus nouis utens legibus et caerimoniis insuper et regis scita contemnens et optime nosti quod non expediat regno tuo ut insolescat per licentiam 3:9 si tibi pla- 3:9 If it please the king, let it be written that they may be

destroyed: and I will pay ten thousand talents of silver to the hands of those that have the charge of the business, to

cet decerne ut pereat et decem milia talentorum adpendam bring it into the king’s treasuries. arcariis gazae tuae 3:10 tulit ergo rex anulum quo utebatur 3:10 And the king took his ring from his hand, and gave

it unto Haman the son of Hammedatha the Agagite, the Jews’ enemy.

de manu sua et dedit eum Aman filio Amadathi de progenie Agag hosti Iudaeorum 3:11 dixitque ad eum argentum 3:11 And the king said unto Haman, The silver is given to

thee, the people also, to do with them as it seemeth good to thee.

quod polliceris tuum sit de populo age quod tibi placet 3:12

3:12 Then were the king’s scribes called on the thirteenth day of the first month, and there was written according to all that Haman had commanded unto the king’s lieutenants, and to the governors that were over every province, and to the rulers of every people of every province according to the writing thereof, and to every people after their language; in the name of king Ahasuerus was it written, and sealed with the king’s ring.

uocatique sunt scribae regis mense primo nisan tertiadecima die eius et scriptum est ut iusserat Aman ad omnes satra-

Biblia Sacra Vulgata

pas regis et iudices prouinciarum diuersarumque gentium ut quaeque gens legere poterat et audire pro uarietate linguarum ex nomine regis Asueri et litterae ipsius signatae anulo

3:13 And the letters were sent by posts into all the king’s 3:13 missae sunt per cursores regis ad uniuersas prouincias

provinces, to destroy, to kill, and to cause to perish, all Jews, both young and old, little children and women, in one day, even upon the thirteenth day of the twelfth month, which is the month Adar, and to take the spoil of them for a prey.

ut occiderent atque delerent omnes Iudaeos a puero usque ad senem paruulos et mulieres uno die hoc est tertiodecimo mensis duodecimi qui uocatur adar et bona eorum diriperent

3:14 The copy of the writing for a commandment to be 3:14 summa autem epistularum haec fuit ut omnes prouinciae

given in every province was published unto all people, that they should be ready against that day.

scirent et pararent se ad praedictam diem 3:15 festinabant

3:15 The posts went out, being hastened by the king’s commandment, and the decree was given in Shushan the palace. And the king and Haman sat down to drink; but the city Shushan was perplexed.

cursores qui missi erant explere regis imperium statimque in Susis pependit edictum rege et Aman celebrante conuiuium et cunctis qui in urbe erant flentibus

4:1 When Mordecai perceived all that was done, Mordecai 4:1 quae cum audisset Mardocheus scidit uestimenta sua rent his clothes, and put on sackcloth with ashes, and went out into the midst of the city, and cried with a loud and a

bitter cry; et indutus est sacco spargens cinerem capiti et in platea me-

diae ciuitatis uoce magna clamabat ostendens amaritudinem

4:2 And came even before the king’s gate: for none might animi sui 4:2 et hoc heiulatu usque ad fores palatii gradiens

enter into the king’s gate clothed with sackcloth.

non enim erat licitum indutum sacco aulam regis intrare

4:3 And in every province, whithersoever the king’s com- 4:3 in omnibus quoque prouinciis oppidis ac locis ad quae mandment and his decree came, there was great mourning among the Jews, and fasting, and weeping, and wailing;

and many lay in sackcloth and ashes. crudele regis dogma peruenerat planctus ingens erat apud

Iudaeos ieiunium ululatus et fletus sacco et cinere multis pro

4:4 So Esther’s maids and her chamberlains came and told strato utentibus 4:4 ingressae sunt autem puellae Hester et

it her. Then was the queen exceedingly grieved; and she sent raiment to clothe Mordecai, and to take away his sack-

cloth from him: but he received it not. eunuchi nuntiaueruntque ei quod audiens consternata est et

misit uestem ut ablato sacco induerent eum quam accipere

4:5 Then called Esther for Hatach, one of the king’s cham- noluit 4:5 accitoque Athac eunucho quem rex ministrum ei berlains, whom he had appointed to attend upon her, and gave him a commandment to Mordecai, to know what it

was, and why it was. dederat praecepit ut iret ad Mardocheum et disceret ab eo 4:6 So Hatach went forth to Mordecai unto the street of cur hoc faceret 4:6 egressusque Athac iuit ad Mardocheum

the city, which was before the king’s gate.

4:7 And Mordecai told him of all that had happened stantem in platea ciuitatis ante ostium palatii 4:7 qui indi-

unto him, and of the sum of the money that Haman had promised to pay to the king’s treasuries for the Jews, to

destroy them. cauit ei omnia quae acciderant quomodo Aman promisisset

ut in thesauros regis pro Iudaeorum nece inferret argentum

4:8 Also he gave him the copy of the writing of the decree 4:8 exemplarque edicti quod pendebat in Susis dedit ei ut

that was given at Shushan to destroy them, to shew it unto Esther, and to declare it unto her, and to charge her that she should go in unto the king, to make supplication unto him, and to make request before him for her people.

reginae ostenderet et moneret eam ut intraret ad regem et

4:9 And Hatach came and told Esther the words of Morde- deprecaretur eum pro populo suo 4:9 regressus Athac nun-

cai.

4:10 Again Esther spake unto Hatach, and gave him com- tiauit Hester omnia quae Mardocheus dixerat 4:10 quae re-

mandment unto Mordecai;

spondit ei et iussit ut diceret Mardocheo 4:11 omnes serui 4:11 All the king’s servants, and the people of the king’s

provinces, do know, that whosoever, whether man or women, shall come unto the king into the inner court, who is not called, there is one law of his to put him to death, except such to whom the king shall hold out the golden sceptre, that he may live: but I have not been called to come in unto the king these thirty days.

regis et cunctae quae sub dicione eius sunt norunt prouinciae quod siue uir siue mulier inuocatus interius atrium regis intrauerit absque ulla cunctatione statim interficiatur nisi forte rex auream uirgam ad eum tetenderit pro signo clementiae atque ita possit uiuere ego igitur quomodo ad regem intrare potero quae triginta iam diebus non sum uocata ad eum 4:12 quod cum audisset Mardocheus 4:13 rursum manda- 4:12 And they told to Mordecai Esther’s words.

4:13 Then Mordecai commanded to answer Esther, Think not with thyself that thou shalt escape in the king’s house, more than all the Jews.

uit Hester dicens ne putes quod animam tuam tantum libe-

res quia in domo regis es prae cunctis Iudaeis 4:14 si enim 4:14 For if thou altogether holdest thy peace at this time, then shall there enlargement and deliverance arise to the Jews from another place; but thou and thy father’s house shall be destroyed: and who knoweth whether thou art come to the kingdom for such a time as this?

nunc silueris per aliam occasionem liberabuntur Iudaei et tu et domus patris tui peribitis et quis nouit utrum idcirco ad regnum ueneris ut in tali tempore parareris 4:15 rursumque 4:15 Then Esther bade them return Mordecai this answer,

Hester haec Mardocheo uerba mandauit 4:16 uade et con- 4:16 Go, gather together all the Jews that are present in Shushan, and fast ye for me, and neither eat nor drink three days, night or day: I also and my maidens will fast likewise; and so will I go in unto the king, which is not according to the law: and if I perish, I perish.

grega omnes Iudaeos quos in Susis reppereris et orate pro me non comedatis et non bibatis tribus diebus ac noctibus et ego cum ancillulis meis similiter ieiunabo et tunc ingrediar ad regem contra legem faciens inuocata tradensque me morti et periculo 4:17 iuit itaque Mardocheus et fecit omnia quae 4:17 So Mordecai went his way, and did according to all

that Esther had commanded him.

ei Hester praeceperat

5:1 die autem tertio induta est Hester regalibus uestimentis 5:1 Now it came to pass on the third day, that Esther put

on her royal apparel, and stood in the inner court of the king’s house, over against the king’s house: and the king sat upon his royal throne in the royal house, over against the gate of the house.

et stetit in atrio domus regiae quod erat interius contra basilicam regis at ille sedebat super solium in consistorio palatii contra ostium domus 5:2 cumque uidisset Hester reginam 5:2 And it was so, when the king saw Esther the queen

standing in the court, that she obtained favour in his sight: and the king held out to Esther the golden sceptre that was in his hand. So Esther drew near, and touched the top of the sceptre.

stantem placuit oculis eius et extendit contra eam uirgam auream quam tenebat manu quae accedens osculata est summitatem uirgae eius 5:3 dixitque ad eam rex quid uis Hester 5:3 Then said the king unto her, What wilt thou, queen

Esther? and what is thy request? it shall be even given thee to the half of the kingdom.

regina quae est petitio tua etiam si dimidiam regni partem petieris dabitur tibi 5:4 at illa respondit si regi placet ob- 5:4 And Esther answered, If it seem good unto the king,

let the king and Haman come this day unto the banquet that I have prepared for him.

secro ut uenias ad me hodie et Aman tecum ad conuiuium quod paraui 5:5 statimque rex uocate inquit cito Aman ut 5:5 Then the king said, Cause Haman to make haste, that

he may do as Esther hath said. So the king and Haman came to the banquet that Esther had prepared.

Hester oboediat uoluntati uenerunt itaque rex et Aman ad conuiuium quod eis regina parauerat 5:6 dixitque ei rex 5:6 And the king said unto Esther at the banquet of wine,

What is thy petition? and it shall be granted thee: and what is thy request? even to the half of the kingdom it shall be performed.

Biblia Sacra Vulgata

postquam uinum biberat abundanter quid petis ut detur tibi et pro qua re postulas etiam si dimidiam partem regni mei

5:7 Then answered Esther, and said, My petition and my petieris inpetrabis 5:7 cui respondit Hester petitio mea et

request is;

5:8 If I have found favour in the sight of the king, and if preces istae sunt 5:8 si inueni gratiam in conspectu regis et

it please the king to grant my petition, and to perform my request, let the king and Haman come to the banquet that I shall prepare for them, and I will do to morrow as the king hath said.

si regi placet ut det mihi quod postulo et meam impleat petitionem ueniat rex et Aman ad conuiuium quod paraui eis et

5:9 Then went Haman forth that day joyful and with a glad cras regi aperiam uoluntatem meam 5:9 egressus est itaque

heart: but when Haman saw Mordecai in the king’s gate, that he stood not up, nor moved for him, he was full of

indignation against Mordecai. illo die Aman laetus et alacer cumque uidisset Mardocheum sedentem ante fores palatii et non solum non adsurrexisse sibi sed nec motum quidem de loco sessionis suae indig-

5:10 Nevertheless Haman refrained himself: and when he natus est ualde 5:10 et dissimulata ira reuersus in domum

came home, he sent and called for his friends, and Zeresh his wife.

suam conuocauit ad se amicos et Zares uxorem suam 5:11

5:11 And Haman told them of the glory of his riches, and the multitude of his children, and all the things wherein the king had promoted him, and how he had advanced him above the princes and servants of the king.

et exposuit illis magnitudinem diuitiarum suarum filiorumque turbam et quanta eum gloria super omnes principes et

5:12 Haman said moreover, Yea, Esther the queen did let seruos suos rex eleuasset 5:12 et post haec ait regina quo-

no man come in with the king unto the banquet that she had prepared but myself; and to morrow am I invited unto

her also with the king. que Hester nullum alium uocauit cum rege ad conuiuium

praeter me apud quam etiam cras cum rege pransurus sum

5:13 Yet all this availeth me nothing, so long as I see 5:13 et cum haec omnia habeam nihil me habere puto quam-

Mordecai the Jew sitting at the king’s gate.

diu uidero Mardocheum Iudaeum sedentem ante fores regias

5:14 Then said Zeresh his wife and all his friends unto 5:14 responderuntque ei Zares uxor eius et ceteri amici iube

him, Let a gallows be made of fifty cubits high, and to morrow speak thou unto the king that Mordecai may be hanged thereon: then go thou in merrily with the king unto the banquet. And the thing pleased Haman; and he caused the gallows to be made.

parari excelsam trabem habentem altitudinem quinquaginta cubitos et dic mane regi ut adpendatur super eam Mardocheus et sic ibis cum rege laetus ad conuiuium placuit ei consilium et iussit excelsam parari crucem

6:1 On that night could not the king sleep, and he com- 6:1 noctem illam rex duxit insomnem iussitque adferri sibi

manded to bring the book of records of the chronicles; and they were read before the king.

historias et annales priorum temporum qui cum illo prae-

6:2 And it was found written, that Mordecai had told of sente legerentur 6:2 uentum est ad eum locum ubi scrip-

Bigthana and Teresh, two of the king’s chamberlains, the keepers of the door, who sought to lay hand on the king

Ahasuerus. tum erat quomodo nuntiasset Mardocheus insidias Bagathan

et Thares eunuchorum regem Asuerum iugulare cupientium

6:3 And the king said, What honour and dignity hath been 6:3 quod cum rex audisset ait quid pro hac fide honoris ac done to Mordecai for this? Then said the king’s servants that ministered unto him, There is nothing done for him.

praemii Mardocheus consecutus est dixeruntque ei serui il-

6:4 And the king said, Who is in the court? Now Haman lius ac ministri nihil omnino mercedis accepit 6:4 statim-

was come into the outward court of the king’s house, to speak unto the king to hang Mordecai on the gallows that

he had prepared for him. que rex quis est inquit in atrio Aman quippe interius atrium domus regiae intrauerat ut suggereret regi et iuberet Mardocheum adfigi patibulo quod ei fuerat praeparatum 6:5 res- 6:5 And the king’s servants said unto him, Behold, Haman

standeth in the court. And the king said, Let him come in.

ponderunt pueri Aman stat in atrio dixitque rex ingrediatur 6:6 cumque esset ingressus ait illi quid debet fieri uiro quem 6:6 So Haman came in. And the king said unto him,

What shall be done unto the man whom the king delighteth to honour? Now Haman thought in his heart, To whom

rex honorare desiderat cogitans Aman in corde suo et re- would the king delight to do honour more than to myself? putans quod nullum alium rex nisi se uellet honorare 6:7 6:7 And Haman answered the king, For the man whom the

king delighteth to honour,

respondit homo quem rex honorare cupit 6:8 debet indui 6:8 Let the royal apparel be brought which the king useth to wear, and the horse that the king rideth upon, and the crown royal which is set upon his head:

uestibus regiis et inponi super equum qui de sella regis est

et accipere regium diadema super caput suum 6:9 et pri- 6:9 And let this apparel and horse be delivered to the hand of one of the king’s most noble princes, that they may array the man withal whom the king delighteth to honour, and bring him on horseback through the street of the city, and proclaim before him, Thus shall it be done to the man whom the king delighteth to honour.

mus de regis principibus ac tyrannis teneat equum eius et per plateam ciuitatis incedens clamet ac dicat sic honorabitur quemcumque rex uoluerit honorare 6:10 dixitque ei rex 6:10 Then the king said to Haman, Make haste, and take

the apparel and the horse, as thou hast said, and do even so to Mordecai the Jew, that sitteth at the king’s gate: let

festina et sumpta stola et equo fac ita ut locutus es Mardo- nothing fail of all that thou hast spoken. cheo Iudaeo qui sedet ante fores palatii caue ne quicquam de his quae locutus es praetermittas 6:11 tulit itaque Aman 6:11 Then took Haman the apparel and the horse, and ar-

rayed Mordecai, and brought him on horseback through the street of the city, and proclaimed before him, Thus shall it be done unto the man whom the king delighteth to honour.

stolam et equum indutumque Mardocheum in platea ciuitatis et inpositum equo praecedebat atque clamabat hoc honore condignus est quemcumque rex uoluerit honorare 6:12 re- 6:12 And Mordecai came again to the king’s gate. But

Haman hasted to his house mourning, and having his head covered.

uersus est Mardocheus ad ianuam palatii et Aman festinauit ire in domum suam lugens et operto capite 6:13 narrauit- 6:13 And Haman told Zeresh his wife and all his friends

every thing that had befallen him. Then said his wise men and Zeresh his wife unto him, If Mordecai be of the seed of the Jews, before whom thou hast begun to fall, thou shalt not prevail against him, but shalt surely fall before him.

que Zares uxori suae et amicis omnia quae euenissent sibi cui responderunt sapientes quos habebat in consilio et uxor eius si de semine Iudaeorum est Mardocheus ante quem cadere coepisti non poteris ei resistere sed cades in conspectu eius 6:14 adhuc illis loquentibus uenerunt eunuchi regis et 6:14 And while they were yet talking with him, came the

king’s chamberlains, and hasted to bring Haman unto the banquet that Esther had prepared.

cito eum ad conuiuium quod regina parauerat pergere conpulerunt

7:1 intrauit itaque rex et Aman ut biberent cum regina 7:2 7:1 So the king and Haman came to banquet with Esther

the queen.

7:2 And the king said again unto Esther on the second day at the banquet of wine, What is thy petition, queen Esther? and it shall be granted thee: and what is thy request? and it shall be performed, even to the half of the kingdom.

dixitque ei rex etiam in secundo die postquam uino incaluerat quae est petitio tua Hester ut detur tibi et quid uis fieri etiam si dimidiam regni mei partem petieris inpetrabis 7:3 7:3 Then Esther the queen answered and said, If I have

found favour in thy sight, O king, and if it please the king, let my life be given me at my petition, and my people at

ad quem illa respondit si inueni gratiam in oculis tuis o rex my request: et si tibi placet dona mihi animam meam pro qua rogo et po-

Biblia Sacra Vulgata

7:4 For we are sold, I and my people, to be destroyed, to pulum meum pro quo obsecro 7:4 traditi enim sumus ego be slain, and to perish. But if we had been sold for bondmen and bondwomen, I had held my tongue, although the

enemy could not countervail the king’s damage. et populus meus ut conteramur iugulemur et pereamus atque utinam in seruos et famulas uenderemur esset tolerabile malum et gemens tacerem nunc autem hostis noster est cuius

7:5 Then the king Ahasuerus answered and said unto Es- crudelitas redundat in regem 7:5 respondensque rex Asue-

ther the queen, Who is he, and where is he, that durst presume in his heart to do so?

rus ait quis est iste et cuius potentiae ut haec audeat facere

7:6 And Esther said, The adversary and enemy is this 7:6 dixit Hester hostis et inimicus noster pessimus iste est

wicked Haman. Then Haman was afraid before the king and the queen.

Aman quod ille audiens ilico obstipuit uultum regis ac regi-

7:7 And the king arising from the banquet of wine in his nae ferre non sustinens 7:7 rex autem surrexit iratus et de

wrath went into the palace garden: and Haman stood up to make request for his life to Esther the queen; for he saw

that there was evil determined against him by the king. loco conuiuii intrauit in hortum arboribus consitum Aman

quoque surrexit ut rogaret Hester reginam pro anima sua in-

7:8 Then the king returned out of the palace garden into tellexit enim a rege sibi paratum malum 7:8 qui cum rethe place of the banquet of wine; and Haman was fallen upon the bed whereon Esther was. Then said the king, Will he force the queen also before me in the house? As the word went out of king’s mouth, they covered Haman’s face.

uersus esset de horto nemoribus consito et intrasset conuiuii locum repperit Aman super lectulum corruisse in quo iacebat Hester et ait etiam reginam uult opprimere me praesente in domo mea necdum uerbum de ore regis exierat et statim

7:9 And Harbonah, one of the chamberlains, said before operuerunt faciem eius 7:9 dixitque Arbona unus de eunu-

the king, Behold also, the gallows fifty cubits high, which Haman had made for Mordecai, who spoken good for the king, standeth in the house of Haman. Then the king said, Hang him thereon.

chis qui stabant in ministerio regis en lignum quod parauerat Mardocheo qui locutus est pro rege stat in domo Aman habens altitudinis quinquaginta cubitos cui dixit rex adpendite

7:10 So they hanged Haman on the gallows that he had eum in eo 7:10 suspensus est itaque Aman in patibulo quod

prepared for Mordecai. Then was the king’s wrath pacified.

parauerat Mardocheo et regis ira quieuit

8:1 On that day did the king Ahasuerus give the house 8:1 die illo dedit rex Asuerus Hester reginae domum of Haman the Jews’ enemy unto Esther the queen. And Mordecai came before the king; for Esther had told what

he was unto her. Aman aduersarii Iudaeorum et Mardocheus ingressus est

ante faciem regis confessa est enim ei Hester quod esset pa-

8:2 And the king took off his ring, which he had taken truus suus 8:2 tulitque rex anulum quem ab Aman recipi

from Haman, and gave it unto Mordecai. And Esther set Mordecai over the house of Haman.

iusserat et tradidit Mardocheo Hester autem constituit Mar-

8:3 And Esther spake yet again before the king, and fell docheum super domum suam 8:3 nec his contenta procidit down at his feet, and besought him with tears to put away the mischief of Haman the Agagite, and his device that he

had devised against the Jews. ad pedes regis fleuitque et locuta ad eum orauit ut malitiam

Aman Agagitae et machinationes eius pessimas quas exco-

8:4 Then the king held out the golden sceptre toward Es- gitauerat contra Iudaeos iuberet irritas fieri 8:4 at ille ex

ther. So Esther arose, and stood before the king,

more sceptrum aureum protendit manu quo signum clemen-

8:5 And said, If it please the king, and if I have favour in tiae monstrabatur illaque consurgens stetit ante eum 8:5 et

his sight, and the thing seem right before the king, and I be pleasing in his eyes, let it be written to reverse the letters devised by Haman the son of Hammedatha the Agagite, which he wrote to destroy the Jews which are in all the king’s provinces:

DCCCL

ait si placet regi et inueni gratiam coram oculis eius et deprecatio mea non ei uidetur esse contraria obsecro ut nouis epistulis ueteres Aman litterae insidiatoris et hostis Iudaeorum quibus eos in cunctis regis prouinciis perire praeceperat corrigantur 8:6 quomodo enim potero sustinere necem 8:6 For how can I endure to see the evil that shall come

unto my people? or how can I endure to see the destruction of my kindred?

et interfectionem populi mei 8:7 responditque rex Asuerus

8:7 Then the king Ahasuerus said unto Esther the queen and to Mordecai the Jew, Behold, I have given Esther the house of Haman, and him they have hanged upon the gallows, because he laid his hand upon the Jews.

Hester reginae et Mardocheo Iudaeo domum Aman concessi Hester et ipsum iussi adfigi cruci qui ausus est manum in Iudaeos mittere 8:8 scribite ergo Iudaeis sicut uobis placet ex 8:8 Write ye also for the Jews, as it liketh you, in the

king’s name, and seal it with the king’s ring: for the writing which is written in the king’s name, and sealed with

regis nomine signantes litteras anulo meo haec enim con- the king’s ring, may no man reverse. suetudo erat ut epistulis quae ex regis nomine mittebantur et illius anulo signatae erant nemo auderet contradicere 8:9 8:9 Then were the king’s scribes called at that time in the

third month, that is, the month Sivan, on the three and twentieth day thereof; and it was written according to all that Mordecai commanded unto the Jews, and to the lieutenants, and the deputies and rulers of the provinces which are from India unto Ethiopia, an hundred twenty and seven provinces, unto every province according to the writing thereof, and unto every people after their language, and to the Jews according to their writing, and according to their language.

accitisque scribis et librariis regis erat autem tempus tertii mensis qui appellatur siban uicesima et tertia illius die scriptae sunt epistulae ut Mardocheus uoluerat ad Iudaeos et ad principes procuratoresque et iudices qui centum uiginti septem prouinciis ab India usque Aethiopiam praesidebant prouinciae atque prouinciae populo et populo iuxta linguas et litteras suas et Iudaeis ut legere poterant et audire 8:10 8:10 And he wrote in the king Ahasuerus’ name, and

sealed it with the king’s ring, and sent letters by posts on horseback, and riders on mules, camels, and young

ipsaeque epistulae quae ex regis nomine mittebantur anulo dromedaries: illius obsignatae sunt et missae per ueredarios qui per omnes prouincias discurrentes ueteres litteras nouis nuntiis praeuenirent 8:11 quibus imperauit rex ut conuenirent Iudaeos per 8:11 Wherein the king granted the Jews which were in

every city to gather themselves together, and to stand for their life, to destroy, to slay and to cause to perish, all the power of the people and province that would assault them, both little ones and women, and to take the spoil of them for a prey,

singulas ciuitates et in unum praeciperent congregari ut starent pro animabus suis et omnes inimicos suos cum coniugibus ac liberis et uniuersis domibus interficerent atque delerent 8:12 et constituta est per omnes prouincias una ul- 8:12 Upon one day in all the provinces of king Ahasuerus,

namely, upon the thirteenth day of the twelfth month, which is the month Adar.

tionis dies id est tertiadecima mensis duodecimi adar 8:13

8:13 The copy of the writing for a commandment to be given in every province was published unto all people, and that the Jews should be ready against that day to avenge themselves on their enemies.

summaque epistulae fuit ut in omnibus terris ac populis qui regis Asueri imperio subiacebant notum fieret paratos esse Iudaeos ad capiendam uindictam de hostibus suis 8:14 eg- 8:14 So the posts that rode upon mules and camels went

out, being hastened and pressed on by the king’s commandment. And the decree was given at Shushan the

ressique sunt ueredarii celeres nuntios perferentes et edictum palace. regis pependit in Susis 8:15 Mardocheus autem de palatio 8:15 And Mordecai went out from the presence of the king

in royal apparel of blue and white, and with a great crown of gold, and with a garment of fine linen and purple: and the city of Shushan rejoiced and was glad.

Biblia Sacra Vulgata

et de conspectu regis egrediens fulgebat uestibus regiis hyacinthinis uidelicet et aerinis coronam auream portans capite et amictus pallio serico atque purpureo omnisque ciuitas ex-

8:16 The Jews had light, and gladness, and joy, and hon- ultauit atque laetata est 8:16 Iudaeis autem noua lux oriri

our. 8:17 And in every province, and in every city, whitherso- uisa est gaudium honor et tripudium 8:17 apud omnes poever the king’s commandment and his decree came, the Jews had joy and gladness, a feast and a good day. And many of the people of the land became Jews; for the fear of the Jews fell upon them.

pulos urbes atque prouincias quocumque regis iussa ueniebant mira exultatio epulae atque conuiuia et festus dies in tantum ut plures alterius gentis et sectae eorum religioni et caerimoniis iungerentur grandis enim cunctos iudaici nominis terror inuaserat

9:1 Now in the twelfth month, that is, the month Adar, on 9:1 igitur duodecimi mensis quem adar uocari ante iam the thirteenth day of the same, when the king’s commandment and his decree drew near to be put in execution, in the day that the enemies of the Jews hoped to have power over them, (though it was turned to the contrary, that the Jews had rule over them that hated them;)

diximus tertiadecima die quando cunctis Iudaeis interfectio parabatur et hostes eorum inhiabant sanguini uersa uice Iudaei superiores esse coeperunt et se de aduersariis uindicare

9:2 The Jews gathered themselves together in their cities 9:2 congregatique sunt per singulas ciuitates oppida et loca

throughout all the provinces of the king Ahasuerus, to lay hand on such as sought their hurt: and no man could with-

stand them; for the fear of them fell upon all people. ut extenderent manum contra inimicos et persecutores suos

nullusque ausus est resistere eo quod omnes populos magni-

9:3 And all the rulers of the provinces, and the lieutenants, tudinis eorum formido penetrarat 9:3 nam et prouinciarum and the deputies, and officers of the king, helped the Jews; because the fear of Mordecai fell upon them.

iudices duces et procuratores omnisque dignitas quae singulis locis et operibus praeerat extollebant Iudaeos timore

9:4 For Mordecai was great in the king’s house, and his Mardochei 9:4 quem principem esse palatii et plurimum

fame went out throughout all the provinces: for this man Mordecai waxed greater and greater.

posse cognouerant fama quoque nominis eius crescebat co-

9:5 Thus the Jews smote all their enemies with the stroke tidie et per cunctorum ora uolitabat 9:5 itaque percusserunt

of the sword, and slaughter, and destruction, and did what they would unto those that hated them.

Iudaei inimicos suos plaga magna et occiderunt eos redden-

9:6 And in Shushan the palace the Jews slew and destroyed tes eis quod sibi parauerant facere 9:6 in tantum ut etiam

five hundred men.

in Susis quingentos uiros interficerent et decem extra filios Aman Agagitae hostis Iudaeorum quorum ista sunt nomina

9:7 And Parshandatha, and Dalphon, and Aspatha, 9:7 Pharsandatha et Delphon et Esphata 9:8 et Phorata et

9:8 And Poratha, and Adalia, and Aridatha, 9:9 And Parmashta, and Arisai, and Aridai, and Vajezatha, Adalia et Aridatha 9:9 et Ephermesta et Arisai et Aridai et

9:10 The ten sons of Haman the son of Hammedatha, the Vaizatha 9:10 quos cum occidissent praedas de substantiis

enemy of the Jews, slew they; but on the spoil laid they not their hand.

eorum agere noluerunt 9:11 statimque numerus eorum qui

9:11 On that day the number of those that were slain in Shushan the palace was brought before the king.

9:12 And the king said unto Esther the queen, The Jews occisi erant in Susis ad regem relatus est 9:12 qui dixit regi-

have slain and destroyed five hundred men in Shushan the palace, and the ten sons of Haman; what have they done in the rest of the king’s provinces? now what is thy petition? and it shall be granted thee: or what is thy request further? and it shall be done.

nae in urbe Susis interfecere Iudaei quingentos uiros et alios decem filios Aman quantam putas eos exercere caedem in uniuersis prouinciis quid ultra postulas et quid uis ut fieri iubeam 9:13 cui illa respondit si regi placet detur potestas 9:13 Then said Esther, If it please the king, let it be granted

to the Jews which are in Shushan to do to morrow also according unto this day’s decree, and let Haman’s ten sons

Iudaeis ut sicut hodie fecerunt in Susis sic et cras faciant be hanged upon the gallows. et decem filii Aman in patibulis suspendantur 9:14 praece- 9:14 And the king commanded it so to be done: and the

decree was given at Shushan; and they hanged Haman’s ten sons.

pitque rex ut ita fieret statimque in Susis pependit edictum et decem Aman filii suspensi sunt 9:15 congregatis Iudaeis 9:15 For the Jews that were in Shushan gathered them-

selves together on the fourteenth day also of the month Adar, and slew three hundred men at Shushan; but on the

quartadecima adar mensis die interfecti sunt in Susis trecenti prey they laid not their hand. uiri nec eorum ab illis direpta substantia est 9:16 sed et per 9:16 But the other Jews that were in the king’s provinces

gathered themselves together, and stood for their lives, and had rest from their enemies, and slew of their foes seventy

omnes prouincias quae dicioni regis subiacebant pro anima- and five thousand, but they laid not their hands on the prey, bus suis stetere Iudaei interfectis hostibus ac persecutoribus suis in tantum ut septuaginta quinque milia occisorum implerentur et nullus de substantiis eorum quicquam contingeret 9:17 dies autem tertiusdecimus mensis adar unus apud 9:17 On the thirteenth day of the month Adar; and on the

fourteenth day of the same rested they, and made it a day of feasting and gladness.

omnes interfectionis fuit et quartodecimo die caedere desierunt quem constituerunt esse sollemnem ut in eo omni deinceps tempore uacarent epulis gaudio atque conuiuiis 9:18 9:18 But the Jews that were at Shushan assembled to-

gether on the thirteenth day thereof, and on the fourteenth thereof; and on the fifteenth day of the same they rested,

at hii qui in urbe Susis caedem exercuerant tertiodecimo et and made it a day of feasting and gladness. quartodecimo eiusdem mensis die in caede uersati sunt quintodecimo autem die percutere desierunt et idcirco eandem diem constituere sollemnem epularum atque laetitiae 9:19 9:19 Therefore the Jews of the villages, that dwelt in the

unwalled towns, made the fourteenth day of the month Adar a day of gladness and feasting, and a good day, and

hii uero Iudaei qui in oppidis non muratis ac uillis mora- of sending portions one to another. bantur quartumdecimum diem mensis adar conuiuiorum et gaudii decreuerunt ita ut exultent in eo et mittant sibi mutuo partes epularum et ciborum 9:20 scripsit itaque Mar- 9:20 And Mordecai wrote these things, and sent letters

unto all the Jews that were in all the provinces of the king Ahasuerus, both nigh and far,

docheus omnia haec et litteris conprehensa misit ad Iudaeos qui in omnibus regis prouinciis morabantur tam in uicino positis quam procul 9:21 ut quartamdecimam et quintamdeci- 9:21 To stablish this among them, that they should keep

the fourteenth day of the month Adar, and the fifteenth day of the same, yearly,

mam diem mensis adar pro festis susciperent et reuertente semper anno sollemni honore celebrarent 9:22 quia in ipsis 9:22 As the days wherein the Jews rested from their en-

emies, and the month which was turned unto them from sorrow to joy, and from mourning into a good day: that they should make them days of feasting and joy, and of sending portions one to another, and gifts to the poor.

diebus se ulti sunt Iudaei de inimicis suis et luctus atque tristitia in hilaritatem gaudiumque conuersa sint essentque istae dies epularum atque laetitiae et mitterent sibi inuicem cibo-

Biblia Sacra Vulgata

9:23 And the Jews undertook to do as they had begun, and rum partes et pauperibus munuscula largirentur 9:23 susce-

as Mordecai had written unto them;

peruntque Iudaei in sollemnem ritum cuncta quae eo tempore facere coeperant et quae Mardocheus litteris facienda

9:24 Because Haman the son of Hammedatha, the mandauerat 9:24 Aman enim filius Amadathi stirpis Agag

Agagite, the enemy of all the Jews, had devised against the Jews to destroy them, and had cast Pur, that is, the lot,

to consume them, and to destroy them; hostis et aduersarius Iudaeorum cogitauit contra eos malum

ut occideret illos atque deleret et misit phur quod nostra lin-

9:25 But when Esther came before the king, he com- gua uertitur in sortem 9:25 et postea ingressa est Hester ad

manded by letters that his wicked device, which he devised against the Jews, should return upon his own head,

and that he and his sons should be hanged on the gallows. regem obsecrans ut conatus eius litteris regis irriti fierent et

malum quod contra Iudaeos cogitauerat reuerteretur in caput

9:26 Wherefore they called these days Purim after the eius denique et ipsum et filios eius adfixerunt cruci 9:26

name of Pur. Therefore for all the words of this letter, and of that which they had seen concerning this matter,

and which had come unto them, atque ex illo tempore dies isti appellati sunt Phurim id est Sortium eo quod phur id est sors in urnam missa fuerit et cuncta quae gesta sunt epistulae id est libri huius uolumine

9:27 The Jews ordained, and took upon them, and upon continentur 9:27 quaeque sustinuerint et quae deinceps in-

their seed, and upon all such as joined themselves unto them, so as it should not fail, that they would keep these two days according to their writing, and according to their appointed time every year;

mutata sint suscepere Iudaei super se et semen suum et super cunctos qui religioni eorum uoluerint copulari ut nulli liceat duos hos dies absque sollemnitate transigere quam scriptura testatur et certa expetunt tempora annis sibi iugiter

9:28 And that these days should be remembered and kept succedentibus 9:28 isti sunt dies quos nulla umquam dele-

throughout every generation, every family, every province, and every city; and that these days of Purim should not fail from among the Jews, nor the memorial of them perish from their seed.

bit obliuio et per singulas generationes cunctae in toto orbe prouinciae celebrabunt nec est ulla ciuitas in qua dies Phurim id est Sortium non obseruentur a Iudaeis et ab eorum

9:29 Then Esther the queen, the daughter of Abihail, and progenie quae his caerimoniis obligata est 9:29 scripserunt-

Mordecai the Jew, wrote with all authority, to confirm this second letter of Purim.

que Hester regina filia Abiahil et Mardocheus Iudaeus etiam secundam epistulam ut omni studio dies ista sollemnis san-

9:30 And he sent the letters unto all the Jews, to the hun- ciretur in posterum 9:30 et miserunt ad omnes Iudaeos qui

dred twenty and seven provinces of the kingdom of Ahasuerus, with words of peace and truth,

in centum uiginti septem regis Asueri prouinciis uersabantur

9:31 To confirm these days of Purim in their times ap- ut haberent pacem et susciperent ueritatem 9:31 obseruan-

pointed, according as Mordecai the Jew and Esther the queen had enjoined them, and as they had decreed for themselves and for their seed, the matters of the fastings and their cry.

tes dies Sortium et suo tempore cum gaudio celebrarent sicut constituerat Mardocheus et Hester et illi obseruanda susceperant a se et a semine suo ieiunia atque clamores et Sortium

9:32 And the decree of Esther confirmed these matters of dies 9:32 et omnia quae libri huius qui uocatur Hester his-

Purim; and it was written in the book.

toria continentur

10:1 rex uero Asuerus omnem terram et cunctas maris in- 10:1 And the king Ahasuerus laid a tribute upon the land, and upon the isles of the sea.

sulas fecit tributarias 10:2 cuius fortitudo et imperium et 10:2 And all the acts of his power and of his might, and

the declaration of the greatness of Mordecai, whereunto the king advanced him, are they not written in the book of

dignitas atque sublimitas qua exaltauit Mardocheum scripta the chronicles of the kings of Media and Persia?

sunt in libris Medorum atque Persarum 10:3 et quomodo 10:3 For Mordecai the Jew was next unto king Ahasuerus, and great among the Jews, and accepted of the multitude of his brethren, seeking the wealth of his people, and

Mardocheus iudaici generis secundus a rege Asuero fuerit et speaking peace to all his seed. magnus inter Iudaeos et acceptabilis plebi fratrum suorum quaerens bona populo suo et loquens ea quae ad pacem sui seminis pertinerent

Biblia Sacra Vulgata

DCCCLVI

\underline{Liber Iob}

1:1 uir erat in terra Hus nomine Iob et erat uir ille simplex 1:1 There was a man in the land of Uz, whose name was

Job; and that man was perfect and upright, and one that feared God, and eschewed evil.

et rectus ac timens Deum et recedens a malo 1:2 natique

1:2 And there were born unto him seven sons and three daughters.

sunt ei septem filii et tres filiae 1:3 et fuit possessio eius se- _{1:3 His substance also was seven thousand sheep, and}

three thousand camels, and five hundred yoke of oxen, and five hundred she asses, and a very great household; so that

ptem milia ouium et tria milia camelorum quingenta quoque _{this man was the greatest of all the men of the east.} iuga boum et quingentae asinae ac familia multa nimis eratque uir ille magnus inter omnes Orientales 1:4 et ibant filii 1:4 And his sons went and feasted in their houses, every

one his day; and sent and called for their three sisters to eat and to drink with them.

eius et faciebant conuiuium per domos unusquisque in die suo et mittentes uocabant tres sorores suas ut comederent et biberent cum eis 1:5 cumque in orbem transissent dies con- 1:5 And it was so, when the days of their feasting were

gone about, that Job sent and sanctified them, and rose up early in the morning, and offered burnt offerings according to the number of them all: for Job said, It may be that my sons have sinned, and cursed God in their hearts. Thus did Job continually.

uiuii mittebat ad eos Iob et sanctificabat illos consurgensque diluculo offerebat holocausta per singulos dicebat enim ne forte peccauerint filii mei et benedixerint Deo in cordibus suis sic faciebat Iob cunctis diebus 1:6 quadam autem 1:6 Now there was a day when the sons of God came to

present themselves before the LORD, and Satan came also among them.

die cum uenissent filii Dei ut adsisterent coram Domino adfuit inter eos etiam Satan 1:7 cui dixit Dominus unde ue- 1:7 And the LORD said unto Satan, Whence comest thou?

Then Satan answered the LORD, and said, From going to and fro in the earth, and from walking up and down in it.

nis qui respondens ait circuiui terram et perambulaui eam 1:8 dixitque Dominus ad eum numquid considerasti seruum 1:8 And the LORD said unto Satan, Hast thou considered

my servant Job, that there is none like him in the earth, a perfect and an upright man, one that feareth God, and

meum Iob quod non sit ei similis in terra homo simplex et escheweth evil? rectus et timens Deum ac recedens a malo 1:9 cui respon- 1:9 Then Satan answered the LORD, and said, Doth Job

fear God for nought?

dens Satan ait numquid frustra timet Iob Deum 1:10 nonne 1:10 Hast not thou made an hedge about him, and about

his house, and about all that he hath on every side? thou hast blessed the work of his hands, and his substance is

tu uallasti eum ac domum eius uniuersamque substantiam increased in the land. per circuitum operibus manuum eius benedixisti et possessio illius creuit in terra 1:11 sed extende paululum manum 1:11 But put forth thine hand now, and touch all that he

hath, and he will curse thee to thy face.

tuam et tange cuncta quae possidet nisi in facie tua benedixerit tibi 1:12 dixit ergo Dominus ad Satan ecce uniuersa 1:12 And the LORD said unto Satan, Behold, all that he

hath is in thy power; only upon himself put not forth thine hand. So Satan went forth from the presence of the LORD.

quae habet in manu tua sunt tantum in eum ne extendas manum tuam egressusque est Satan a facie Domini 1:13 cum 1:13 And there was a day when his sons and his daugh-

ters were eating and drinking wine in their eldest brother’s house:

autem quadam die filii et filiae eius comederent et biberent uinum in domo fratris sui primogeniti 1:14 nuntius uenit ad 1:14 And there came a messenger unto Job, and said, The

oxen were plowing, and the asses feeding beside them:

Iob qui diceret boues arabant et asinae pascebantur iuxta eos 1:15 et inruerunt Sabei tuleruntque omnia et pueros percus- 1:15 And the Sabeans fell upon them, and took them away;

yea, they have slain the servants with the edge of the sword; and I only am escaped alone to tell thee.

serunt gladio et euasi ego solus ut nuntiarem tibi 1:16 cum-

1:16 While he was yet speaking, there came also another, and said, The fire of God is fallen from heaven, and hath burned up the sheep, and the servants, and consumed them; and I only am escaped alone to tell thee.

que adhuc ille loqueretur uenit alter et dixit ignis Dei cecidit

Biblia Sacra Vulgata

e caelo et tactas oues puerosque consumpsit et effugi ego

1:17 While he was yet speaking, there came also another, solus ut nuntiarem tibi 1:17 sed et illo adhuc loquente uenit

and said, The Chaldeans made out three bands, and fell upon the camels, and have carried them away, yea, and slain the servants with the edge of the sword; and I only am escaped alone to tell thee.

alius et dixit Chaldei fecerunt tres turmas et inuaserunt camelos et tulerunt eos necnon et pueros percusserunt gladio

1:18 While he was yet speaking, there came also another, et ego fugi solus ut nuntiarem tibi 1:18 loquebatur ille et

and said, Thy sons and thy daughters were eating and drinking wine in their eldest brother’s house:

ecce alius intrauit et dixit filiis tuis et filiabus uescentibus et

1:19 And, behold, there came a great wind from the bibentibus uinum in domo fratris sui primogeniti 1:19 re-

wilderness, and smote the four corners of the house, and it fell upon the young men, and they are dead; and I only am

escaped alone to tell thee. pente uentus uehemens inruit a regione deserti et concussit

quattuor angulos domus quae corruens oppressit liberos tuos

1:20 Then Job arose, and rent his mantle, and shaved his et mortui sunt et effugi ego solus ut nuntiarem tibi 1:20 tunc

head, and fell down upon the ground, and worshipped,

surrexit Iob et scidit tunicam suam et tonso capite corruens

1:21 And said, Naked came I out of my mother’s womb, in terram adorauit 1:21 et dixit nudus egressus sum de utero

and naked shall I return thither: the LORD gave, and the LORD hath taken away; blessed be the name of the

LORD. matris meae et nudus reuertar illuc Dominus dedit Dominus

1:22 In all this Job sinned not, nor charged God foolishly. abstulit sit nomen Domini benedictum 1:22 in omnibus his

non peccauit Iob neque stultum quid contra Deum locutus est

2:1 Again there was a day when the sons of God came to 2:1 factum est autem cum quadam die uenissent filii Dei et

present themselves before the LORD, and Satan came also among them to present himself before the LORD.

starent coram Domino uenisset quoque Satan inter eos et sta-

2:2 And the LORD said unto Satan, From whence comest ret in conspectu eius 2:2 ut diceret Dominus ad Satan unde

thou? And Satan answered the LORD, and said, From going to and fro in the earth, and from walking up and

down in it. uenis qui respondens ait circuiui terram et perambulaui eam 2:3 And the LORD said unto Satan, Hast thou considered 2:3 et dixit Dominus ad Satan numquid considerasti seruum

my servant Job, that there is none like him in the earth, a perfect and an upright man, one that feareth God, and escheweth evil? and still he holdeth fast his integrity, although thou movedst me against him, to destroy him without cause.

meum Iob quod non sit ei similis in terra uir simplex et rectus timens Deum ac recedens a malo et adhuc retinens innocentiam tu autem commouisti me aduersus eum ut adfligerem

2:4 And Satan answered the LORD, and said, Skin for illum frustra 2:4 cui respondens Satan ait pellem pro pelle

skin, yea, all that a man hath will he give for his life.

2:5 But put forth thine hand now, and touch his bone and et cuncta quae habet homo dabit pro anima sua 2:5 alioquin

his flesh, and he will curse thee to thy face.

mitte manum tuam et tange os eius et carnem et tunc uide-

2:6 And the LORD said unto Satan, Behold, he is in thine bis quod in facie benedicat tibi 2:6 dixit ergo Dominus ad

hand; but save his life.

Satan ecce in manu tua est uerumtamen animam illius serua

2:7 So went Satan forth from the presence of the LORD, 2:7 egressus igitur Satan a facie Domini percussit Iob ulcere

and smote Job with sore boils from the sole of his foot unto his crown.

pessimo a planta pedis usque ad uerticem eius 2:8 qui testa

2:8 And he took him a potsherd to scrape himself withal; and he sat down among the ashes.

2:9 Then said his wife unto him, Dost thou still retain thine saniem deradebat sedens in sterquilinio 2:9 dixit autem illi

integrity? curse God, and die.

uxor sua adhuc tu permanes in simplicitate tua benedic Deo et morere 2:10 qui ait ad illam quasi una de stultis locuta 2:10 But he said unto her, Thou speakest as one of the

foolish women speaketh. What? shall we receive good at the hand of God, and shall we not receive evil? In all this

es si bona suscepimus de manu Domini quare mala non su- did not Job sin with his lips. scipiamus in omnibus his non peccauit Iob labiis suis 2:11 2:11 Now when Job’s three friends heard of all this evil

that was come upon him, they came every one from his own place; Eliphaz the Temanite, and Bildad the Shuhite, and Zophar the Naamathite: for they had made an appointment together to come to mourn with him and to comfort him.

igitur audientes tres amici Iob omne malum quod accidisset ei uenerunt singuli de loco suo Eliphaz Themanites et Baldad Suites et Sophar Naamathites condixerant enim ut pariter uenientes uisitarent eum et consolarentur 2:12 cumque 2:12 And when they lifted up their eyes afar off, and knew

him not, they lifted up their voice, and wept; and they rent every one his mantle, and sprinkled dust upon their heads

leuassent procul oculos suos non cognouerunt eum et excla- toward heaven. mantes plorauerunt scissisque uestibus sparserunt puluerem super caput suum in caelum 2:13 et sederunt cum eo in ter- 2:13 So they sat down with him upon the ground seven

days and seven nights, and none spake a word unto him: for they saw that his grief was very great.

ram septem diebus et septem noctibus et nemo loquebatur ei uerbum uidebant enim dolorem esse uehementem

3:1 post haec aperuit Iob os suum et maledixit diei suo 3:1 After this opened Job his mouth, and cursed his day.

3:2 et locutus est 3:3 pereat dies in qua natus sum et nox in 3:2 And Job spake, and said,

3:3 Let the day perish wherein I was born, and the night in which it was said, There is a man child conceived.

qua dictum est conceptus est homo 3:4 dies ille uertatur in

3:4 Let that day be darkness; let not God regard it from above, neither let the light shine upon it.

tenebras non requirat eum Deus desuper et non inlustret lumine 3:5 obscurent eum tenebrae et umbra mortis occupet 3:5 Let darkness and the shadow of death stain it; let a

cloud dwell upon it; let the blackness of the day terrify it.

eum caligo et inuoluatur amaritudine 3:6 noctem illam te- 3:6 As for that night, let darkness seize upon it; let it not

be joined unto the days of the year, let it not come into the number of the months.

nebrosus turbo possideat non conputetur in diebus anni nec numeretur in mensibus 3:7 sit nox illa solitaria nec laude 3:7 Lo, let that night be solitary, let no joyful voice come

therein.

digna 3:8 maledicant ei qui maledicunt diei qui parati sunt 3:8 Let them curse it that curse the day, who are ready to

raise up their mourning.

suscitare Leuiathan 3:9 obtenebrentur stellae caligine eius 3:9 Let the stars of the twilight thereof be dark; let it look

for light, but have none; neither let it see the dawning of the day:

expectet lucem et non uideat nec ortum surgentis aurorae 3:10 quia non conclusit ostia uentris qui portauit me nec ab- 3:10 Because it shut not up the doors of my mother’s

womb, nor hid sorrow from mine eyes.

stulit mala ab oculis meis 3:11 quare non in uulua mortuus 3:11 Why died I not from the womb? why did I not give

up the ghost when I came out of the belly?

sum egressus ex utero non statim perii 3:12 quare exceptus 3:12 Why did the knees prevent me? or why the breasts

that I should suck?

genibus cur lactatus uberibus 3:13 nunc enim dormiens si- 3:13 For now should I have lain still and been quiet, I

should have slept: then had I been at rest,

lerem et somno meo requiescerem 3:14 cum regibus et con- 3:14 With kings and counsellors of the earth, which build

desolate places for themselves;

sulibus terrae qui aedificant sibi solitudines 3:15 aut cum 3:15 Or with princes that had gold, who filled their houses

with silver:

principibus qui possident aurum et replent domos suas argento 3:16 aut sicut abortiuum absconditum non subsiste- 3:16 Or as an hidden untimely birth I had not been; as

infants which never saw light.

rem uel qui concepti non uiderunt lucem 3:17 ibi impii ces- 3:17 There the wicked cease from troubling; and there the

weary be at rest.

sauerunt a tumultu et ibi requieuerunt fessi robore 3:18 et 3:18 There the prisoners rest together; they hear not the

voice of the oppressor.

quondam uincti pariter sine molestia non audierunt uocem

3:19 The small and great are there; and the servant is free exactoris 3:19 paruus et magnus ibi sunt et seruus liber a from his master. 3:20 Wherefore is light given to him that is in misery, and domino suo 3:20 quare data est misero lux et uita his qui life unto the bitter in soul;

3:21 Which long for death, but it cometh not; and dig for in amaritudine animae sunt 3:21 qui expectant mortem et

it more than for hid treasures;

3:22 Which rejoice exceedingly, and are glad, when they non uenit quasi effodientes thesaurum 3:22 gaudentque ue-

can find the grave?

3:23 Why is light given to a man whose way is hid, and hementer cum inuenerint sepulchrum 3:23 uiro cuius abs-

whom God hath hedged in?

3:24 For my sighing cometh before I eat, and my roarings condita est uia et circumdedit eum Deus tenebris 3:24 ante-

are poured out like the waters.

quam comedam suspiro et quasi inundantes aquae sic rugitus

3:25 For the thing which I greatly feared is come upon me, meus 3:25 quia timor quem timebam euenit mihi et quod and that which I was afraid of is come unto me.

3:26 I was not in safety, neither had I rest, neither was I uerebar accidit 3:26 nonne dissimulaui nonne silui nonne

quiet; yet trouble came.

quieui et uenit super me indignatio

4:1 Then Eliphaz the Temanite answered and said, 4:1 respondens autem Eliphaz Themanites dixit 4:2 si

4:2 If we assay to commune with thee, wilt thou be grieved? but who can withhold himself from speaking?

coeperimus loqui tibi forsitan moleste accipias sed conce-

4:3 Behold, thou hast instructed many, and thou hast ptum sermonem tenere quis possit 4:3 ecce docuisti mul-

strengthened the weak hands.

4:4 Thy words have upholden him that was falling, and tos et manus lassas roborasti 4:4 uacillantes confirmaue-

thou hast strengthened the feeble knees.

4:5 But now it is come upon thee, and thou faintest; it runt sermones tui et genua trementia confortasti 4:5 nunc

toucheth thee, and thou art troubled.

autem uenit super te plaga et defecisti tetigit te et conturba-

4:6 Is not this thy fear, thy confidence, thy hope, and the tus es 4:6 timor tuus fortitudo tua patientia tua et perfec-

uprightness of thy ways? 4:7 Remember, I pray thee, who ever perished, being in- tio uiarum tuarum 4:7 recordare obsecro te quis umquam nocent? or where were the righteous cut off?

4:8 Even as I have seen, they that plow iniquity, and sow innocens perierit aut quando recti deleti sint 4:8 quin po-

wickedness, reap the same.

tius uidi eos qui operantur iniquitatem et seminant dolores

4:9 By the blast of God they perish, and by the breath of et metunt eos 4:9 flante Deo perisse et spiritu irae eius esse

his nostrils are they consumed.

4:10 The roaring of the lion, and the voice of the fierce consumptos 4:10 rugitus leonis et uox leaenae et dentes ca-

lion, and the teeth of the young lions, are broken.

4:11 The old lion perisheth for lack of prey, and the stout tulorum leonum contriti sunt 4:11 tigris periit eo quod non

lion’s whelps are scattered abroad.

4:12 Now a thing was secretly brought to me, and mine haberet praedam et catuli leonis dissipati sunt 4:12 porro

ear received a little thereof.

ad me dictum est uerbum absconditum et quasi furtiue sus-

4:13 In thoughts from the visions of the night, when deep cepit auris mea uenas susurri eius 4:13 in horrore uisionis

sleep falleth on men,

4:14 Fear came upon me, and trembling, which made all nocturnae quando solet sopor occupare homines 4:14 pauor

my bones to shake.

4:15 Then a spirit passed before my face; the hair of my tenuit me et tremor et omnia ossa mea perterrita sunt 4:15

flesh stood up:

et cum spiritus me praesente transiret inhorruerunt pili car-

4:16 It stood still, but I could not discern the form thereof: nis meae 4:16 stetit quidam cuius non agnoscebam uultum an image was before mine eyes, there was silence, and I heard a voice, saying,

imago coram oculis meis et uocem quasi aurae lenis audiui

4:17 Shall mortal man be more just than God? shall a man 4:17 numquid homo Dei conparatione iustificabitur aut fac-

be more pure than his maker? tore suo purior erit uir 4:18 ecce qui seruiunt ei non sunt 4:18 Behold, he put no trust in his servants; and his angels

he charged with folly:

stabiles et in angelis suis repperit prauitatem 4:19 quanto 4:19 How much less in them that dwell in houses of clay,

whose foundation is in the dust, which are crushed before the moth?

magis hii qui habitant domos luteas qui terrenum habent fundamentum consumentur uelut a tinea 4:20 de mane usque 4:20 They are destroyed from morning to evening: they

perish for ever without any regarding it.

ad uesperum succidentur et quia nullus intellegit in aeternum peribunt 4:21 qui autem reliqui fuerint auferentur ex 4:21 Doth not their excellency which is in them go away?

they die, even without wisdom.

eis morientur et non in sapientia

5:1 uoca ergo si est qui tibi respondeat et ad aliquem 5:1 Call now, if there be any that will answer thee; and to

which of the saints wilt thou turn?

sanctorum conuertere 5:2 uere stultum interficit iracundia 5:2 For wrath killeth the foolish man, and envy slayeth the

silly one.

et paruulum occidit inuidia 5:3 ego uidi stultum firma ra- 5:3 I have seen the foolish taking root: but suddenly I

cursed his habitation.

dice et maledixi pulchritudini eius statim 5:4 longe fient 5:4 His children are far from safety, and they are crushed

in the gate, neither is there any to deliver them.

filii eius a salute et conterentur in porta et non erit qui eruat 5:5 cuius messem famelicus comedet et ipsum rapiet arma- 5:5 Whose harvest the hungry eateth up, and taketh it even

out of the thorns, and the robber swalloweth up their substance.

tus et ebibent sitientes diuitias eius 5:6 nihil in terra sine

5:6 Although affliction cometh not forth of the dust, neither doth trouble spring out of the ground;

causa fit et de humo non orietur dolor 5:7 homo ad laborem 5:7 Yet man is born unto trouble, as the sparks fly upward.

nascitur et auis ad uolatum 5:8 quam ob rem ego depre- 5:8 I would seek unto God, and unto God would I commit

my cause:

cabor Dominum et ad Deum ponam eloquium meum 5:9 5:9 Which doeth great things and unsearchable; marvel-

lous things without number:

qui facit magna et inscrutabilia et mirabilia absque numero 5:10 qui dat pluuiam super faciem terrae et inrigat aquis uni- 5:10 Who giveth rain upon the earth, and sendeth waters

upon the fields:

uersa 5:11 qui ponit humiles in sublimi et maerentes eri- 5:11 To set up on high those that be low; that those which

mourn may be exalted to safety.

git sospitate 5:12 qui dissipat cogitationes malignorum ne 5:12 He disappointeth the devices of the crafty, so that

their hands cannot perform their enterprise.

possint implere manus eorum quod coeperant 5:13 qui ad- 5:13 He taketh the wise in their own craftiness: and the

counsel of the froward is carried headlong.

prehendit sapientes in astutia eorum et consilium prauorum dissipat 5:14 per diem incurrent tenebras et quasi in nocte 5:14 They meet with darkness in the day time, and grope

in the noonday as in the night.

sic palpabunt in meridie 5:15 porro saluum faciet a gladio 5:15 But he saveth the poor from the sword, from their

mouth, and from the hand of the mighty.

oris eorum et de manu uiolenti pauperem 5:16 et erit egeno 5:16 So the poor hath hope, and iniquity stoppeth her

mouth.

spes iniquitas autem contrahet os suum 5:17 beatus homo 5:17 Behold, happy is the man whom God correcteth:

therefore despise not thou the chastening of the Almighty:

qui corripitur a Domino increpationem ergo Domini ne reprobes 5:18 quia ipse uulnerat et medetur percutit et manus 5:18 For he maketh sore, and bindeth up: he woundeth,

and his hands make whole.

eius sanabunt 5:19 in sex tribulationibus liberabit te et in se- 5:19 He shall deliver thee in six troubles: yea, in seven

there shall no evil touch thee.

ptima non tanget te malum 5:20 in fame eruet te de morte 5:20 In famine he shall redeem thee from death: and in

war from the power of the sword.

et in bello de manu gladii 5:21 a flagello linguae absconde- 5:21 Thou shalt be hid from the scourge of the tongue:

neither shalt thou be afraid of destruction when it cometh.

ris et non timebis calamitatem cum uenerit 5:22 in uastitate 5:22 At destruction and famine thou shalt laugh: neither

shalt thou be afraid of the beasts of the earth.

5:23 For thou shalt be in league with the stones of the field: et fame ridebis et bestiam terrae non formidabis 5:23 sed and the beasts of the field shall be at peace with thee.

cum lapidibus regionum pactum tuum et bestiae terrae paci-

5:24 And thou shalt know that thy tabernacle shall be in ficae erunt tibi 5:24 et scies quod pacem habeat tabernacu-

peace; and thou shalt visit thy habitation, and shalt not sin.

lum tuum et uisitans speciem tuam non peccabis 5:25 scies

5:25 Thou shalt know also that thy seed shall be great, and thine offspring as the grass of the earth.

quoque quoniam multiplex erit semen tuum et progenies tua

5:26 Thou shalt come to thy grave in a full age, like as a quasi herba terrae 5:26 ingredieris in abundantia sepulch-

shock of corn cometh in in his season. 5:27 Lo this, we have searched it, so it is; hear it, and know rum sicut infertur aceruus in tempore suo 5:27 ecce hoc ut thou it for thy good.

inuestigauimus ita est quod auditum mente pertracta

6:1 But Job answered and said, 6:1 respondens autem Iob dixit 6:2 utinam adpenderentur

6:2 Oh that my grief were throughly weighed, and my calamity laid in the balances together!

peccata mea quibus iram merui et calamitas quam patior in

6:3 For now it would be heavier than the sand of the sea: statera 6:3 quasi harena maris haec grauior appareret unde

therefore my words are swallowed up. 6:4 For the arrows of the Almighty are within me, the poi- et uerba mea dolore sunt plena 6:4 quia sagittae Domini son whereof drinketh up my spirit: the terrors of God do set themselves in array against me.

in me sunt quarum indignatio ebibit spiritum meum et ter-

6:5 Doth the wild ass bray when he hath grass? or loweth rores Domini militant contra me 6:5 numquid rugiet ona-

the ox over his fodder?

ger cum habuerit herbam aut mugiet bos cum ante praesepe

6:6 Can that which is unsavoury be eaten without salt? or plenum steterit 6:6 aut poterit comedi insulsum quod non

is there any taste in the white of an egg?

est sale conditum aut potest aliquis gustare quod gustatum

6:7 The things that my soul refused to touch are as my adfert mortem 6:7 quae prius tangere nolebat anima mea

sorrowful meat.

6:8 Oh that I might have my request; and that God would nunc prae angustia cibi mei sunt 6:8 quis det ut ueniat peti-

grant me the thing that I long for!

6:9 Even that it would please God to destroy me; that he tio mea et quod expecto tribuat mihi Deus 6:9 et qui coepit

would let loose his hand, and cut me off! 6:10 Then should I yet have comfort; yea, I would harden ipse me conterat soluat manum suam et succidat me 6:10 myself in sorrow: let him not spare; for I have not concealed the words of the Holy One.

et haec mihi sit consolatio ut adfligens me dolore non parcat

6:11 What is my strength, that I should hope? and what is nec contradicam sermonibus Sancti 6:11 quae est enim for-

mine end, that I should prolong my life?

titudo mea ut sustineam aut quis finis meus ut patienter agam

6:12 Is my strength the strength of stones? or is my flesh 6:12 nec fortitudo lapidum fortitudo mea nec caro mea aerea

of brass?

6:13 Is not my help in me? and is wisdom driven quite est 6:13 ecce non est auxilium mihi in me et necessarii quo-

from me?

6:14 To him that is afflicted pity should be shewed from que mei recesserunt a me 6:14 qui tollit ab amico suo mi-

his friend; but he forsaketh the fear of the Almighty.

6:15 My brethren have dealt deceitfully as a brook, and as sericordiam timorem Domini derelinquit 6:15 fratres mei

the stream of brooks they pass away;

praeterierunt me sicut torrens qui raptim transit in conualli-

6:16 Which are blackish by reason of the ice, and wherein bus 6:16 qui timent pruinam inruet super eos nix 6:17 tem-

the snow is hid: 6:17 What time they wax warm, they vanish: when it is hot, they are consumed out of their place.

pore quo fuerint dissipati peribunt et ut incaluerit soluentur

6:18 The paths of their way are turned aside; they go to de loco suo 6:18 inuolutae sunt semitae gressuum eorum

nothing, and perish.

6:19 The troops of Tema looked, the companies of Sheba ambulabunt in uacuum et peribunt 6:19 considerate semi-

waited for them.

tas Theman itinera Saba et expectate paulisper 6:20 confusi 6:20 They were confounded because they had hoped; they

came thither, and were ashamed.

sunt quia speraui uenerunt quoque usque ad me et pudore cooperti sunt 6:21 nunc uenistis et modo uidentes plagam 6:21 For now ye are nothing; ye see my casting down, and

are afraid.

meam timetis 6:22 numquid dixi adferte mihi et de substan- 6:22 Did I say, Bring unto me? or, Give a reward for me

of your substance?

tia uestra donate mihi 6:23 uel liberate me de manu hostis 6:23 Or, Deliver me from the enemy’s hand? or, Redeem

me from the hand of the mighty?

et de manu robustorum eruite me 6:24 docete me et ego 6:24 Teach me, and I will hold my tongue: and cause me

to understand wherein I have erred.

tacebo et si quid forte ignoraui instruite me 6:25 quare de- 6:25 How forcible are right words! but what doth your

arguing reprove?

traxistis sermonibus ueritatis cum e uobis nullus sit qui possit arguere 6:26 ad increpandum tantum eloquia concinna- 6:26 Do ye imagine to reprove words, and the speeches of

one that is desperate, which are as wind?

tis et in uentum uerba profertis 6:27 super pupillum inruitis 6:27 Yea, ye overwhelm the fatherless, and ye dig a pit for

your friend.

et subuertere nitimini amicum uestrum 6:28 uerumtamen 6:28 Now therefore be content, look upon me; for it is

evident unto you if I lie.

quod coepistis explete praebete aurem et uidete an mentiar 6:29 respondete obsecro absque contentione et loquentes id 6:29 Return, I pray you, let it not be iniquity; yea, return

again, my righteousness is in it.

quod iustum est iudicate 6:30 et non inuenietis in lingua 6:30 Is there iniquity in my tongue? cannot my taste dis-

cern perverse things?

mea iniquitatem nec in faucibus meis stultitia personabit

7:1 militia est uita hominis super terram et sicut dies mer- 7:1 Is there not an appointed time to man upon earth? are

not his days also like the days of an hireling?

cennarii dies eius 7:2 sicut seruus desiderat umbram et sic- 7:2 As a servant earnestly desireth the shadow, and as an

hireling looketh for the reward of his work:

ut mercennarius praestolatur finem operis sui 7:3 sic et ego 7:3 So am I made to possess months of vanity, and weari-

some nights are appointed to me.

habui menses uacuos et noctes laboriosas enumeraui mihi 7:4 si dormiero dico quando consurgam et rursum expectabo 7:4 When I lie down, I say, When shall I arise, and the

night be gone? and I am full of tossings to and fro unto the dawning of the day.

uesperam et replebor doloribus usque ad tenebras 7:5 in-

7:5 My flesh is clothed with worms and clods of dust; my skin is broken, and become loathsome.

duta est caro mea putredine et sordibus pulueris cutis mea aruit et contracta est 7:6 dies mei uelocius transierunt quam 7:6 My days are swifter than a weaver’s shuttle, and are

spent without hope.

a texente tela succiditur et consumpti sunt absque ulla spe 7:7 memento quia uentus est uita mea et non reuertetur ocu- 7:7 O remember that my life is wind: mine eye shall no

more see good.

lus meus ut uideat bona 7:8 nec aspiciet me uisus hominis 7:8 The eye of him that hath seen me shall see me no more:

thine eyes are upon me, and I am not.

oculi tui in me et non subsistam 7:9 sicut consumitur nu- 7:9 As the cloud is consumed and vanisheth away: so he

that goeth down to the grave shall come up no more.

bes et pertransit sic qui descenderit ad inferos non ascendet

7:10 nec reuertetur ultra in domum suam neque cognoscet 7:10 He shall return no more to his house, neither shall his

place know him any more.

eum amplius locus eius 7:11 quapropter et ego non parcam 7:11 Therefore I will not refrain my mouth; I will speak in the anguish of my spirit; I will complain in the bitterness of my soul.

ori meo loquar in tribulatione spiritus mei confabulabor cum

amaritudine animae meae 7:12 numquid mare sum ego aut 7:12 Am I a sea, or a whale, that thou settest a watch over

me?

cetus quia circumdedisti me carcere 7:13 si dixero conso- 7:13 When I say, My bed shall comfort me, my couch shall

ease my complaints;

labitur me lectulus meus et releuabor loquens mecum in st-

7:14 Then thou scarest me with dreams, and terrifiest me rato meo 7:14 terrebis me per somnia et per uisiones hor-

through visions: 7:15 So that my soul chooseth strangling, and death rather rore concuties 7:15 quam ob rem elegit suspendium anima than my life. 7:16 I loathe it; I would not live alway: let me alone; for mea et mortem ossa mea 7:16 desperaui nequaquam ultra my days are vanity. 7:17 What is man, that thou shouldest magnify him? and iam uiuam parce mihi nihil enim sunt dies mei 7:17 quid that thou shouldest set thine heart upon him?

est homo quia magnificas eum aut quia ponis erga eum cor

7:18 And that thou shouldest visit him every morning, and tuum 7:18 uisitas eum diluculo et subito probas illum 7:19

try him every moment? 7:19 How long wilt thou not depart from me, nor let me alone till I swallow down my spittle?

usquequo non parces mihi nec dimittis me ut gluttiam sali-

7:20 I have sinned; what shall I do unto thee, O thou pre- uam meam 7:20 peccaui quid faciam tibi o custos hominum server of men? why hast thou set me as a mark against thee, so that I am a burden to myself?

quare posuisti me contrarium tibi et factus sum mihimet ipsi

7:21 And why dost thou not pardon my transgression, and grauis 7:21 cur non tolles peccatum meum et quare non au-

take away my iniquity? for now shall I sleep in the dust; and thou shalt seek me in the morning, but I shall not be.

feres iniquitatem meam ecce nunc in puluere dormiam et si mane me quaesieris non subsistam

8:1 Then answered Bildad the Shuhite, and said, 8:1 respondens autem Baldad Suites dixit 8:2 usquequo

8:2 How long wilt thou speak these things? and how long shall the words of thy mouth be like a strong wind?

loqueris talia et spiritus multiplex sermones oris tui 8:3

8:3 Doth God pervert judgment? or doth the Almighty pervert justice?

numquid Deus subplantat iudicium et Omnipotens subuertit

8:4 If thy children have sinned against him, and he have quod iustum est 8:4 etiam si filii tui peccauerunt ei et di-

cast them away for their transgression;

8:5 If thou wouldest seek unto God betimes, and make thy misit eos in manu iniquitatis suae 8:5 tu tamen si diluculo

supplication to the Almighty;

consurrexeris ad Deum et Omnipotentem fueris deprecatus

8:6 If thou wert pure and upright; surely now he would 8:6 si mundus et rectus incesseris statim euigilabit ad te et

awake for thee, and make the habitation of thy righteousness prosperous.

pacatum reddet habitaculum iustitiae tuae 8:7 in tantum ut

8:7 Though thy beginning was small, yet thy latter end should greatly increase.

priora tua fuerint parua et nouissima tua multiplicentur nimis

8:8 For enquire, I pray thee, of the former age, and prepare 8:8 interroga enim generationem pristinam et diligenter in-

thyself to the search of their fathers:

8:9 (For we are but of yesterday, and know nothing, be- uestiga patrum memoriam 8:9 hesterni quippe sumus et ig-

cause our days upon earth are a shadow:)

noramus quoniam sicut umbra dies nostri sunt super terram

8:10 Shall not they teach thee, and tell thee, and utter 8:10 et ipsi docebunt te loquentur tibi et de corde suo pro-

words out of their heart?

8:11 Can the rush grow up without mire? can the flag grow ferent eloquia 8:11 numquid uiuere potest scirpus absque

without water? 8:12 Whilst it is yet in his greenness, and not cut down, it humore aut crescet carectum sine aqua 8:12 cum adhuc sit withereth before any other herb.

8:13 So are the paths of all that forget God; and the hyp- in flore nec carpatur manu ante omnes herbas arescit 8:13

ocrite’s hope shall perish:

sic uiae omnium qui obliuiscuntur Deum et spes hypocritae

8:14 Whose hope shall be cut off, and whose trust shall be peribit 8:14 non ei placebit uecordia sua et sicut tela arane-

a spider’s web. 8:15 He shall lean upon his house, but it shall not stand: arum fiducia eius 8:15 innitetur super domum suam et non he shall hold it fast, but it shall not endure.

8:16 He is green before the sun, and his branch shooteth stabit fulciet eam et non consurget 8:16 humectus uidetur

forth in his garden.

antequam ueniat sol et in horto suo germen eius egreditur 8:17 super aceruum petrarum radices eius densabuntur et in- 8:17 His roots are wrapped about the heap, and seeth the

place of stones.

ter lapides commorabitur 8:18 si absorbuerit eum de loco 8:18 If he destroy him from his place, then it shall deny

him, saying, I have not seen thee.

suo negabit eum et dicet non noui te 8:19 haec est enim lae- 8:19 Behold, this is the joy of his way, and out of the earth

shall others grow.

titia uiae eius ut rursum de terra alii germinentur 8:20 Deus 8:20 Behold, God will not cast away a perfect man, neither

will he help the evil doers:

non proiciet simplicem nec porriget manum malignis 8:21 8:21 Till he fill thy mouth with laughing, and thy lips with

rejoicing.

donec impleatur risu os tuum et labia tua iubilo 8:22 qui 8:22 They that hate thee shall be clothed with shame; and

the dwelling place of the wicked shall come to nought.

oderunt te induentur confusione et tabernaculum impiorum non subsistet

9:1 et respondens Iob ait 9:2 uere scio quod ita sit et quod 9:1 Then Job answered and said,

9:2 I know it is so of a truth: but how should man be just with God?

non iustificetur homo conpositus Deo 9:3 si uoluerit con-

9:3 If he will contend with him, he cannot answer him one of a thousand.

tendere cum eo non poterit ei respondere unum pro mille 9:4 sapiens corde est et fortis robore quis restitit ei et pacem 9:4 He is wise in heart, and mighty in strength: who hath

hardened himself against him, and hath prospered?

habuit 9:5 qui transtulit montes et nescierunt hii quos sub- 9:5 Which removeth the mountains, and they know not:

which overturneth them in his anger.

uertit in furore suo 9:6 qui commouet terram de loco suo 9:6 Which shaketh the earth out of her place, and the pil-

lars thereof tremble.

et columnae eius concutiuntur 9:7 qui praecipit soli et non 9:7 Which commandeth the sun, and it riseth not; and

sealeth up the stars.

oritur et stellas claudit quasi sub signaculo 9:8 qui exten- 9:8 Which alone spreadeth out the heavens, and treadeth

upon the waves of the sea.

dit caelos solus et graditur super fluctus maris 9:9 qui facit 9:9 Which maketh Arcturus, Orion, and Pleiades, and the

chambers of the south.

Arcturum et Oriona et Hyadas et interiora austri 9:10 qui 9:10 Which doeth great things past finding out; yea, and

wonders without number.

facit magna et inconprehensibilia et mirabilia quorum non est numerus 9:11 si uenerit ad me non uidebo si abierit non 9:11 Lo, he goeth by me, and I see him not: he passeth on

also, but I perceive him not.

intellegam eum 9:12 si repente interroget quis respondebit 9:12 Behold, he taketh away, who can hinder him? who

will say unto him, What doest thou?

ei uel quis dicere potest cur facis 9:13 Deus cuius resis- 9:13 If God will not withdraw his anger, the proud helpers

do stoop under him.

tere irae nemo potest et sub quo curuantur qui portant orbem 9:14 quantus ergo sum ego qui respondeam ei et loquar uer- 9:14 How much less shall I answer him, and choose out

my words to reason with him?

bis meis cum eo 9:15 qui etiam si habuero quippiam iustum 9:15 Whom, though I were righteous, yet would I not an-

swer, but I would make supplication to my judge.

non respondebo sed meum iudicem deprecabor 9:16 et cum 9:16 If I had called, and he had answered me; yet would I

not believe that he had hearkened unto my voice.

inuocantem exaudierit me non credo quod audierit uocem meam 9:17 in turbine enim conteret me et multiplicabit uul- 9:17 For he breaketh me with a tempest, and multiplieth

my wounds without cause.

nera mea etiam sine causa 9:18 non concedit requiescere 9:18 He will not suffer me to take my breath, but filleth

me with bitterness.

spiritum meum et implet me amaritudinibus 9:19 si forti- 9:19 If I speak of strength, lo, he is strong: and if of judg-

ment, who shall set me a time to plead?

tudo quaeritur robustissimus est si aequitas iudicii nemo pro me audet testimonium dicere 9:20 si iustificare me uoluero 9:20 If I justify myself, mine own mouth shall condemn

me: if I say, I am perfect, it shall also prove me perverse.

os meum condemnabit me si innocentem ostendere prauum

9:21 Though I were perfect, yet would I not know my soul: me conprobabit 9:21 etiam si simplex fuero hoc ipsum ig-

I would despise my life. 9:22 This is one thing, therefore I said it, He destroyeth norabit anima mea et taedebit me uitae meae 9:22 unum est the perfect and the wicked.

quod locutus sum et innocentem et impium ipse consumit

9:23 If the scourge slay suddenly, he will laugh at the trial 9:23 si flagellat occidat semel et non de poenis innocentum of the innocent.

9:24 The earth is given into the hand of the wicked: he rideat 9:24 terra data est in manu impii uultum iudicum eius

covereth the faces of the judges thereof; if not, where, and who is he?

operit quod si non ille est quis ergo est 9:25 dies mei uelo-

9:25 Now my days are swifter than a post: they flee away, they see no good.

9:26 They are passed away as the swift ships: as the eagle ciores fuerunt cursore fugerunt et non uiderunt bonum 9:26

that hasteth to the prey.

pertransierunt quasi naues poma portantes sicut aquila uo-

9:27 If I say, I will forget my complaint, I will leave off lans ad escam 9:27 cum dixero nequaquam ita loquar commy heaviness, and comfort myself:

9:28 I am afraid of all my sorrows, I know that thou wilt muto faciem meam et dolore torqueor 9:28 uerebar omnia

not hold me innocent.

9:29 If I be wicked, why then labour I in vain? opera mea sciens quod non parceres delinquenti 9:29 si au-9:30 If I wash myself with snow water, and make my tem et sic impius sum quare frustra laboraui 9:30 si lotus

hands never so clean;

fuero quasi aquis niuis et fulserint uelut mundissimae manus

9:31 Yet shalt thou plunge me in the ditch, and mine own meae 9:31 tamen sordibus intingues me et abominabuntur

clothes shall abhor me.

9:32 For he is not a man, as I am, that I should answer me uestimenta mea 9:32 neque enim uiro qui similis mei

him, and we should come together in judgment.

est respondebo nec qui mecum in iudicio ex aequo possit

9:33 Neither is there any daysman betwixt us, that might audiri 9:33 non est qui utrumque ualeat arguere et ponere

lay his hand upon us both. 9:34 Let him take his rod away from me, and let not his manum suam in ambobus 9:34 auferat a me uirgam suam fear terrify me:

9:35 Then would I speak, and not fear him; but it is not so et pauor eius non me terreat 9:35 loquar et non timebo eum

with me.

neque enim possum metuens respondere

10:1 My soul is weary of my life; I will leave my com- 10:1 taedet animam meam uitae meae dimittam aduer-

plaint upon myself; I will speak in the bitterness of my soul.

sum me eloquium meum loquar in amaritudine animae meae

10:2 I will say unto God, Do not condemn me; shew me 10:2 dicam Deo noli me condemnare indica mihi cur me ita

wherefore thou contendest with me. 10:3 Is it good unto thee that thou shouldest oppress, that iudices 10:3 numquid bonum tibi uidetur si calumnieris et thou shouldest despise the work of thine hands, and shine upon the counsel of the wicked?

opprimas me opus manuum tuarum et consilium impiorum

10:4 Hast thou eyes of flesh? or seest thou as man seeth? adiuues 10:4 numquid oculi carnei tibi sunt aut sicut uidet 10:5 Are thy days as the days of man? are thy years as homo et tu uidebis 10:5 numquid sicut dies hominis dies tui

man’s days, 10:6 That thou enquirest after mine iniquity, and searchest et anni tui sicut humana sunt tempora 10:6 ut quaeras inafter my sin?

10:7 Thou knowest that I am not wicked; and there is none iquitatem meam et peccatum meum scruteris 10:7 et scias

that can deliver out of thine hand.

quia nihil impium fecerim cum sit nemo qui de manu tua

10:8 Thine hands have made me and fashioned me to- possit eruere 10:8 manus tuae plasmauerunt me et fecerunt

gether round about; yet thou dost destroy me.

10:9 Remember, I beseech thee, that thou hast made me as me totum in circuitu et sic repente praecipitas me 10:9 me-

the clay; and wilt thou bring me into dust again?

mento quaeso quod sicut lutum feceris me et in puluerem reduces me 10:10 nonne sicut lac mulsisti me et sicut ca- 10:10 Hast thou not poured me out as milk, and curdled

me like cheese?

seum me coagulasti 10:11 pelle et carnibus uestisti me et 10:11 Thou hast clothed me with skin and flesh, and hast

fenced me with bones and sinews.

ossibus et neruis conpegisti me 10:12 uitam et misericor- 10:12 Thou hast granted me life and favour, and thy visi-

tation hath preserved my spirit.

diam tribuisti mihi et uisitatio tua custodiuit spiritum meum 10:13 licet haec celes in corde tuo tamen scio quia uniuerso- 10:13 And these things hast thou hid in thine heart: I know

that this is with thee.

rum memineris 10:14 si peccaui et ad horam pepercisti mihi 10:14 If I sin, then thou markest me, and thou wilt not

acquit me from mine iniquity.

cur ab iniquitate mea mundum me esse non pateris 10:15 et 10:15 If I be wicked, woe unto me; and if I be righteous,

yet will I not lift up my head. I am full of confusion; therefore see thou mine affliction;

si impius fuero uae mihi est et si iustus non leuabo caput saturatus adflictione et miseria 10:16 et propter superbiam 10:16 For it increaseth. Thou huntest me as a fierce lion:

and again thou shewest thyself marvellous upon me.

quasi leaenam capies me reuersusque mirabiliter me crucias 10:17 instauras testes tuos contra me et multiplicas iram tuam 10:17 Thou renewest thy witnesses against me, and in-

creasest thine indignation upon me; changes and war are against me.

aduersum me et poenae militant in me 10:18 quare de uulua

10:18 Wherefore then hast thou brought me forth out of the womb? Oh that I had given up the ghost, and no eye had seen me!

eduxisti me qui utinam consumptus essem ne oculus me uideret 10:19 fuissem quasi qui non essem de utero translatus 10:19 I should have been as though I had not been; I

should have been carried from the womb to the grave.

ad tumulum 10:20 numquid non paucitas dierum meorum 10:20 Are not my days few? cease then, and let me alone,

that I may take comfort a little,

finietur breui dimitte ergo me ut plangam paululum dolorem meum 10:21 antequam uadam et non reuertar ad terram te- 10:21 Before I go whence I shall not return, even to the

land of darkness and the shadow of death;

nebrosam et opertam mortis caligine 10:22 terram miseriae 10:22 A land of darkness, as darkness itself; and of the

shadow of death, without any order, and where the light is as darkness.

et tenebrarum ubi umbra mortis et nullus ordo et sempiternus horror inhabitans

11:1 respondens autem Sophar Naamathites dixit 11:2 11:1 Then answered Zophar the Naamathite, and said,

11:2 Should not the multitude of words be answered? and should a man full of talk be justified?

numquid qui multa loquitur non et audiet aut uir uerbosus iustificabitur 11:3 tibi soli tacebunt homines et cum cete- 11:3 Should thy lies make men hold their peace? and when

thou mockest, shall no man make thee ashamed?

ros inriseris a nullo confutaberis 11:4 dixisti enim purus est 11:4 For thou hast said, My doctrine is pure, and I am

clean in thine eyes.

sermo meus et mundus sum in conspectu tuo 11:5 atque uti- 11:5 But oh that God would speak, and open his lips

against thee;

nam Deus loqueretur tecum et aperiret labia sua tibi 11:6 11:6 And that he would shew thee the secrets of wisdom,

that they are double to that which is! Know therefore that God exacteth of thee less than thine iniquity deserveth.

ut ostenderet tibi secreta sapientiae et quod multiplex esset lex eius et intellegeres quod multo minora exigaris a Deo quam meretur iniquitas tua 11:7 forsitan uestigia Dei con- 11:7 Canst thou by searching find out God? canst thou

find out the Almighty unto perfection?

prehendes et usque ad perfectum Omnipotentem repperies 11:8 excelsior caelo est et quid facies profundior inferno et 11:8 It is as high as heaven; what canst thou do? deeper

than hell; what canst thou know?

unde cognosces 11:9 longior terrae mensura eius et latior 11:9 The measure thereof is longer than the earth, and

broader than the sea.

11:10 If he cut off, and shut up, or gather together, then mari 11:10 si subuerterit omnia uel in unum coartauerit quis

who can hinder him?

11:11 For he knoweth vain men: he seeth wickedness also; contradicet ei 11:11 ipse enim nouit hominum uanitatem et

will he not then consider it?

11:12 For vain men would be wise, though man be born uidens iniquitatem nonne considerat 11:12 uir uanus in su-

like a wild ass’s colt.

perbiam erigitur et tamquam pullum onagri se liberum na-

11:13 If thou prepare thine heart, and stretch out thine tum putat 11:13 tu autem firmasti cor tuum et expandisti ad

hands toward him; 11:14 If iniquity be in thine hand, put it far away, and let eum manus tuas 11:14 si iniquitatem quod est in manu tua not wickedness dwell in thy tabernacles.

abstuleris a te et non manserit in tabernaculo tuo iniustitia

11:15 For then shalt thou lift up thy face without spot; yea, 11:15 tum leuare poteris faciem tuam absque macula et eris thou shalt be stedfast, and shalt not fear:

11:16 Because thou shalt forget thy misery, and remember stabilis et non timebis 11:16 miseriae quoque obliuisceris

it as waters that pass away:

11:17 And thine age shall be clearer than the noonday: et quasi aquarum quae praeterierint recordaberis 11:17 et

thou shalt shine forth, thou shalt be as the morning.

quasi meridianus fulgor consurget tibi ad uesperam et cum

11:18 And thou shalt be secure, because there is hope; yea, te consumptum putaueris orieris ut lucifer 11:18 et habethou shalt dig about thee, and thou shalt take thy rest in safety.

bis fiduciam proposita tibi spe et defossus securus dormies

11:19 Also thou shalt lie down, and none shall make thee 11:19 requiesces et non erit qui te exterreat et deprecabuntur

afraid; yea, many shall make suit unto thee. 11:20 But the eyes of the wicked shall fail, and they shall faciem tuam plurimi 11:20 oculi autem impiorum deficient not escape, and their hope shall be as the giving up of the ghost.

et effugium peribit ab eis et spes eorum abominatio animae

12:1 And Job answered and said, 12:1 respondens autem Iob dixit 12:2 ergo uos estis soli

12:2 No doubt but ye are the people, and wisdom shall die with you.

homines et uobiscum morietur sapientia 12:3 et mihi est cor

12:3 But I have understanding as well as you; I am not inferior to you: yea, who knoweth not such things as these?

sicut et uobis nec inferior uestri sum quis enim haec quae no-

12:4 I am as one mocked of his neighbour, who calleth stis ignorat 12:4 qui deridetur ab amico suo sicut ego inuo-

upon God, and he answereth him: the just upright man is laughed to scorn.

cabit Deum et exaudiet eum deridetur enim iusti simplicitas

12:5 He that is ready to slip with his feet is as a lamp 12:5 lampas contempta apud cogitationes diuitum parata ad

despised in the thought of him that is at ease. 12:6 The tabernacles of robbers prosper, and they that pro- tempus statutum 12:6 abundant tabernacula praedonum et voke God are secure; into whose hand God bringeth abundantly.

audacter prouocant Deum cum ipse dederit omnia in ma-

12:7 But ask now the beasts, and they shall teach thee; and nibus eorum 12:7 nimirum interroga iumenta et docebunt the fowls of the air, and they shall tell thee:

12:8 Or speak to the earth, and it shall teach thee: and the te et uolatilia caeli et indicabunt tibi 12:8 loquere terrae et

fishes of the sea shall declare unto thee.

12:9 Who knoweth not in all these that the hand of the respondebit tibi et narrabunt pisces maris 12:9 quis igno-

LORD hath wrought this? 12:10 In whose hand is the soul of every living thing, and rat quod omnia haec manus Domini fecerit 12:10 in cuius the breath of all mankind.

manu anima omnis uiuentis et spiritus uniuersae carnis ho-

12:11 Doth not the ear try words? and the mouth taste his minis 12:11 nonne auris uerba diiudicat et fauces comeden-

meat?

12:12 With the ancient is wisdom; and in length of days tis saporem 12:12 in antiquis est sapientia et in multo tem-

understanding.

12:13 With him is wisdom and strength, he hath counsel pore prudentia 12:13 apud ipsum est sapientia et fortitudo

and understanding.

12:14 Behold, he breaketh down, and it cannot be built ipse habet consilium et intellegentiam 12:14 si destruxerit

again: he shutteth up a man, and there can be no opening.

nemo est qui aedificet et si incluserit hominem nullus est qui aperiat 12:15 si continuerit aquas omnia siccabuntur et si 12:15 Behold, he withholdeth the waters, and they dry up:

also he sendeth them out, and they overturn the earth.

emiserit eas subuertent terram 12:16 apud ipsum est forti- 12:16 With him is strength and wisdom: the deceived and

the deceiver are his.

tudo et sapientia ipse nouit et decipientem et eum qui decipitur 12:17 adducit consiliarios in stultum finem et iudices in 12:17 He leadeth counsellors away spoiled, and maketh

the judges fools.

stuporem 12:18 balteum regum dissoluit et praecingit fune 12:18 He looseth the bond of kings, and girdeth their loins

with a girdle.

renes eorum 12:19 ducit sacerdotes inglorios et optimates 12:19 He leadeth princes away spoiled, and overthroweth

the mighty.

subplantat 12:20 commutans labium ueracium et doctrinam 12:20 He removeth away the speech of the trusty, and

taketh away the understanding of the aged.

senum auferens 12:21 effundit despectionem super princi- 12:21 He poureth contempt upon princes, and weakeneth

the strength of the mighty.

pes et eos qui oppressi fuerant releuans 12:22 qui reuelat 12:22 He discovereth deep things out of darkness, and

bringeth out to light the shadow of death.

profunda de tenebris et producit in lucem umbram mortis 12:23 qui multiplicat gentes et perdet eas et subuersas in in- 12:23 He increaseth the nations, and destroyeth them: he

enlargeth the nations, and straiteneth them again.

tegrum restituet 12:24 qui inmutat cor principum populi ter- 12:24 He taketh away the heart of the chief of the people

of the earth, and causeth them to wander in a wilderness where there is no way.

rae et decipit eos ut frustra incedant per inuium 12:25 pal-

12:25 They grope in the dark without light, and he maketh them to stagger like a drunken man.

pabunt quasi in tenebris et non in luce et errare eos faciet quasi ebrios

13:1 ecce omnia et uidit oculus meus et audiuit auris mea 13:1 Lo, mine eye hath seen all this, mine ear hath heard

and understood it.

et intellexi singula 13:2 secundum scientiam uestram et ego 13:2 What ye know, the same do I know also: I am not

inferior unto you.

noui nec inferior uestri sum 13:3 sed tamen ad Omnipoten- 13:3 Surely I would speak to the Almighty, and I desire to

reason with God.

tem loquar et disputare cum Deo cupio 13:4 prius uos os- 13:4 But ye are forgers of lies, ye are all physicians of no

value.

tendens fabricatores mendacii et cultores peruersorum dogmatum 13:5 atque utinam taceretis ut putaremini esse sa- 13:5 O that ye would altogether hold your peace! and it

should be your wisdom.

pientes 13:6 audite ergo correptiones meas et iudicium la- 13:6 Hear now my reasoning, and hearken to the pleadings

of my lips.

biorum meorum adtendite 13:7 numquid Deus indiget ue- 13:7 Will ye speak wickedly for God? and talk deceitfully

for him?

stro mendacio ut pro illo loquamini dolos 13:8 numquid fa- 13:8 Will ye accept his person? will ye contend for God? ciem eius accipitis et pro Deo iudicare nitimini 13:9 aut pla- 13:9 Is it good that he should search you out? or as one

man mocketh another, do ye so mock him?

cebit ei quem celare nihil potest aut decipietur ut homo uestris fraudulentiis 13:10 ipse uos arguet quoniam in abscon- 13:10 He will surely reprove you, if ye do secretly accept

persons.

dito faciem eius accipitis 13:11 statim ut se commouerit tur- 13:11 Shall not his excellency make you afraid? and his

dread fall upon you?

babit uos et terror eius inruet super uos 13:12 memoria ues- 13:12 Your remembrances are like unto ashes, your bodies

to bodies of clay.

tra conparabitur cineri et rediientur in lutum ceruices uestrae 13:13 tacete paulisper ut loquar quodcumque mihi mens sug- 13:13 Hold your peace, let me alone, that I may speak, and

let come on me what will.

gesserit 13:14 quare lacero carnes meas dentibus meis et 13:14 Wherefore do I take my flesh in my teeth, and put

my life in mine hand?

13:15 Though he slay me, yet will I trust in him: but I will animam meam porto in manibus meis 13:15 etiam si occimaintain mine own ways before him.

derit me in ipso sperabo uerumtamen uias meas in conspectu

13:16 He also shall be my salvation: for an hypocrite shall eius arguam 13:16 et ipse erit saluator meus non enim uenot come before him. 13:17 Hear diligently my speech, and my declaration with niet in conspectu eius omnis hypocrita 13:17 audite seryour ears.

13:18 Behold now, I have ordered my cause; I know that I monem meum et enigmata percipite auribus uestris 13:18

shall be justified.

13:19 Who is he that will plead with me? for now, if I hold si fuero iudicatus scio quod iustus inueniar 13:19 quis est

my tongue, I shall give up the ghost.

13:20 Only do not two things unto me: then will I not hide qui iudicetur mecum ueniat quare tacens consumor 13:20

myself from thee.

duo tantum ne facias mihi et tunc a facie tua non abscondar

13:21 Withdraw thine hand far from me: and let not thy 13:21 manum tuam longe fac a me et formido tua non me ter-

dread make me afraid.

13:22 Then call thou, and I will answer: or let me speak, reat 13:22 et uoca me et respondebo tibi aut certe loquar et

and answer thou me.

13:23 How many are mine iniquities and sins? make me tu responde mihi 13:23 quantas habeo iniquitates et peccata

to know my transgression and my sin.

13:24 Wherefore hidest thou thy face, and holdest me for scelera mea et delicta ostende mihi 13:24 cur faciem tuam

thine enemy?

13:25 Wilt thou break a leaf driven to and fro? and wilt abscondis et arbitraris me inimicum tuum 13:25 contra fo-

thou pursue the dry stubble?

lium quod uento rapitur ostendis potentiam tuam et stipulam

13:26 For thou writest bitter things against me, and makest siccam persequeris 13:26 scribis enim contra me amaritudi-

me to possess the iniquities of my youth.

13:27 Thou puttest my feet also in the stocks, and lookest nes et consumere me uis peccatis adulescentiae meae 13:27

narrowly unto all my paths; thou settest a print upon the heels of my feet.

posuisti in neruo pedem meum et obseruasti omnes semi-

13:28 And he, as a rotten thing, consumeth, as a garment tas meas et uestigia pedum meorum considerasti 13:28 qui

that is moth eaten.

quasi putredo consumendus sum et quasi uestimentum quod comeditur a tinea

14:1 Man that is born of a woman is of few days and full 14:1 homo natus de muliere breui uiuens tempore reple-

of trouble.

14:2 He cometh forth like a flower, and is cut down: he tus multis miseriis 14:2 quasi flos egreditur et conteritur

fleeth also as a shadow, and continueth not.

et fugit uelut umbra et numquam in eodem statu perma-

14:3 And doth thou open thine eyes upon such an one, and net 14:3 et dignum ducis super huiuscemodi aperire ocu-

bringest me into judgment with thee?

14:4 Who can bring a clean thing out of an unclean? not los tuos et adducere eum tecum in iudicium 14:4 quis po-

one.

test facere mundum de inmundo conceptum semine nonne

14:5 Seeing his days are determined, the number of his tu qui solus es 14:5 breues dies hominis sunt numerus men-

months are with thee, thou hast appointed his bounds that he cannot pass;

suum eius apud te est constituisti terminos eius qui praeterire

14:6 Turn from him, that he may rest, till he shall accom- non poterunt 14:6 recede paululum ab eo ut quiescat doplish, as an hireling, his day.

14:7 For there is hope of a tree, if it be cut down, that it nec optata ueniat sicut mercennarii dies eius 14:7 lignum

will sprout again, and that the tender branch thereof will not cease.

habet spem si praecisum fuerit rursum uirescit et rami eius

14:8 Though the root thereof wax old in the earth, and the pullulant 14:8 si senuerit in terra radix eius et in puluere

stock thereof die in the ground;

14:9 Yet through the scent of water it will bud, and bring emortuus fuerit truncus illius 14:9 ad odorem aquae ger-

forth boughs like a plant.

minabit et faciet comam quasi cum primum plantatum est 14:10 homo uero cum mortuus fuerit et nudatus atque con- 14:10 But man dieth, and wasteth away: yea, man giveth

up the ghost, and where is he?

sumptus ubi quaeso est 14:11 quomodo si recedant aquae 14:11 As the waters fail from the sea, and the flood de-

cayeth and drieth up:

de mari et fluuius uacuefactus arescat 14:12 sic homo cum 14:12 So man lieth down, and riseth not: till the heavens

be no more, they shall not awake, nor be raised out of their sleep.

dormierit non resurget donec adteratur caelum non euigilabit nec consurget de somno suo 14:13 quis mihi hoc tribuat 14:13 O that thou wouldest hide me in the grave, that thou

wouldest keep me secret, until thy wrath be past, that thou wouldest appoint me a set time, and remember me!

ut in inferno protegas me ut abscondas me donec pertranseat furor tuus et constituas mihi tempus in quo recorderis mei 14:14 putasne mortuus homo rursum uiuet cunctis die- 14:14 If a man die, shall he live again? all the days of my

appointed time will I wait, till my change come.

bus quibus nunc milito expecto donec ueniat inmutatio mea 14:15 uocabis et ego respondebo tibi operi manuum tuarum 14:15 Thou shalt call, and I will answer thee: thou wilt

have a desire to the work of thine hands.

porriges dexteram 14:16 tu quidem gressus meos dinume- 14:16 For now thou numberest my steps: dost thou not

watch over my sin?

rasti sed parces peccatis meis 14:17 signasti quasi in sac- 14:17 My transgression is sealed up in a bag, and thou

sewest up mine iniquity.

culo delicta mea sed curasti iniquitatem meam 14:18 mons 14:18 And surely the mountains falling cometh to nought,

and the rock is removed out of his place.

cadens defluet et saxum transfertur de loco suo 14:19 lapi- 14:19 The waters wear the stones: thou washest away the

things which grow out of the dust of the earth; and thou destroyest the hope of man.

des excauant aquae et adluuione paulatim terra consumitur

et homines ergo similiter perdes 14:20 roborasti eum pau- 14:20 Thou prevailest for ever against him, and he passeth:

thou changest his countenance, and sendest him away.

lulum ut in perpetuum pertransiret inmutabis faciem eius et emittes eum 14:21 siue nobiles fuerint filii eius siue ignobi- 14:21 His sons come to honour, and he knoweth it not; and

they are brought low, but he perceiveth it not of them.

les non intelleget 14:22 attamen caro eius dum uiuet dolebit 14:22 But his flesh upon him shall have pain, and his soul

within him shall mourn.

et anima illius super semet ipso lugebit

15:1 respondens autem Eliphaz Themanites dixit 15:2 15:1 Then answered Eliphaz the Temanite, and said,

15:2 Should a wise man utter vain knowledge, and fill his belly with the east wind?

numquid sapiens respondebit quasi in uentum loquens et implebit ardore stomachum suum 15:3 arguis uerbis eum 15:3 Should he reason with unprofitable talk? or with

speeches wherewith he can do no good?

qui non est aequalis tui et loqueris quod tibi non expedit 15:4 quantum in te est euacuasti timorem et tulisti preces co- 15:4 Yea, thou castest off fear, and restrainest prayer be-

fore God.

ram Deo 15:5 docuit enim iniquitas tua os tuum et imitaris 15:5 For thy mouth uttereth thine iniquity, and thou choos-

est the tongue of the crafty.

linguam blasphemantium 15:6 condemnabit te os tuum et 15:6 Thine own mouth condemneth thee, and not I: yea,

thine own lips testify against thee.

non ego et labia tua respondebunt tibi 15:7 numquid pri- 15:7 Art thou the first man that was born? or wast thou

made before the hills?

mus homo tu natus es et ante colles formatus 15:8 num- 15:8 Hast thou heard the secret of God? and dost thou

restrain wisdom to thyself?

quid consilium Dei audisti et inferior te erit eius sapientia 15:9 quid nosti quod ignoremus quid intellegis quod nescia- 15:9 What knowest thou, that we know not? what under-

standest thou, which is not in us?

mus 15:10 et senes et antiqui sunt in nobis multo uetustio- 15:10 With us are both the grayheaded and very aged men,

much elder than thy father.

15:11 Are the consolations of God small with thee? is res quam patres tui 15:11 numquid grande est ut console-

there any secret thing with thee?

15:12 Why doth thine heart carry thee away? and what do tur te Deus sed uerba tua praua hoc prohibent 15:12 quid

thy eyes wink at,

te eleuat cor tuum et quasi magna cogitans adtonitos habes

15:13 That thou turnest thy spirit against God, and lettest oculos 15:13 quid tumet contra Deum spiritus tuus ut prosuch words go out of thy mouth?

15:14 What is man, that he should be clean? and he which feras de ore huiuscemodi sermones 15:14 quid est homo ut

is born of a woman, that he should be righteous?

15:15 Behold, he putteth no trust in his saints; yea, the inmaculatus sit et ut iustus appareat natus de muliere 15:15

heavens are not clean in his sight.

ecce inter sanctos eius nemo inmutabilis et caeli non sunt

15:16 How much more abominable and filthy is man, mundi in conspectu eius 15:16 quanto magis abominabilis

which drinketh iniquity like water?

15:17 I will shew thee, hear me; and that which I have seen et inutilis homo qui bibit quasi aquas iniquitatem 15:17 os-

I will declare;

15:18 Which wise men have told from their fathers, and tendam tibi audi me quod uidi narrabo tibi 15:18 sapientes

have not hid it:

15:19 Unto whom alone the earth was given, and no confitentur et non abscondunt patres suos 15:19 quibus so-

stranger passed among them.

15:20 The wicked man travaileth with pain all his days, lis data est terra et non transibit alienus per eos 15:20 cunc-

and the number of years is hidden to the oppressor.

tis diebus suis impius superbit et numerus annorum incer-

15:21 A dreadful sound is in his ears: in prosperity the tus est tyrannidis eius 15:21 sonitus terroris semper in au-

destroyer shall come upon him.

15:22 He believeth not that he shall return out of darkness, ribus illius et cum pax sit ille insidias suspicatur 15:22 non

and he is waited for of the sword.

credit quod reuerti possit de tenebris circumspectans undi-

15:23 He wandereth abroad for bread, saying, Where is it? que gladium 15:23 cum se mouerit ad quaerendum panem he knoweth that the day of darkness is ready at his hand.

15:24 Trouble and anguish shall make him afraid; they nouit quod paratus sit in manu eius tenebrarum dies 15:24

shall prevail against him, as a king ready to the battle.

terrebit eum tribulatio et angustia uallabit eum sicut regem

15:25 For he stretcheth out his hand against God, and qui praeparatur ad proelium 15:25 tetendit enim aduersus

strengtheneth himself against the Almighty.

Deum manum suam et contra Omnipotentem roboratus est

15:26 He runneth upon him, even on his neck, upon the 15:26 cucurrit aduersus eum erecto collo et pingui ceruice

thick bosses of his bucklers:

15:27 Because he covereth his face with his fatness, and armatus est 15:27 operuit faciem eius crassitudo et de la-

maketh collops of fat on his flanks.

15:28 And he dwelleth in desolate cities, and in houses teribus eius aruina dependet 15:28 habitauit in ciuitatibus

which no man inhabiteth, which are ready to become heaps.

desolatis et in domibus desertis quae in tumulos sunt redac-

15:29 He shall not be rich, neither shall his substance con- tae 15:29 non ditabitur nec perseuerabit substantia eius nec

tinue, neither shall he prolong the perfection thereof upon the earth.

mittet in terra radicem suam 15:30 non recedet de teneb-

15:30 He shall not depart out of darkness; the flame shall dry up his branches, and by the breath of his mouth shall he go away.

ris ramos eius arefaciet flamma et auferetur spiritu oris sui

15:31 Let not him that is deceived trust in vanity: for van- 15:31 non credat frustra errore deceptus quod aliquo pretio

ity shall be his recompence.

15:32 It shall be accomplished before his time, and his rediiendus sit 15:32 antequam dies eius impleantur peribit

branch shall not be green.

15:33 He shall shake off his unripe grape as the vine, and et manus eius arescet 15:33 laedetur quasi uinea in primo

shall cast off his flower as the olive. 15:34 For the congregation of hypocrites shall be desolate, flore botrus eius et quasi oliua proiciens florem suum 15:34 and fire shall consume the tabernacles of bribery.

congregatio enim hypocritae sterilis et ignis deuorabit tabernacula eorum qui munera libenter accipiunt 15:35 concepit 15:35 They conceive mischief, and bring forth vanity, and

their belly prepareth deceit.

dolorem et peperit iniquitatem et uterus eius praeparat dolos

16:1 respondens autem Iob dixit 16:2 audiui frequenter 16:1 Then Job answered and said,

16:2 I have heard many such things: miserable comforters are ye all.

talia consolatores onerosi omnes uos estis 16:3 numquid

16:3 Shall vain words have an end? or what emboldeneth thee that thou answerest?

habebunt finem uerba uentosa aut aliquid tibi molestum est si loquaris 16:4 poteram et ego similia uestri loqui atque 16:4 I also could speak as ye do: if your soul were in my

soul’s stead,

utinam esset anima uestra pro anima mea 16:5 consolarer 16:5 I could heap up words against you, and shake mine

head at you.

et ego uos sermonibus et mouerem caput meum super uos 16:6 roborarem uos ore meo et mouerem labia quasi parcens 16:6 But I would strengthen you with my mouth, and the

moving of my lips should asswage your grief.

uobis 16:7 sed quid agam si locutus fuero non quiescet do- 16:7 Though I speak, my grief is not asswaged: and

though I forbear, what am I eased?

lor meus et si tacuero non recedet a me 16:8 nunc autem 16:8 But now he hath made me weary: thou hast made

desolate all my company.

oppressit me dolor meus et in nihili redacti sunt omnes artus mei 16:9 rugae meae testimonium dicunt contra me et sus- 16:9 And thou hast filled me with wrinkles, which is a wit-

ness against me: and my leanness rising up in me beareth witness to my face.

citatur falsiloquus aduersus faciem meam contradicens mihi 16:10 collegit furorem suum in me et comminans mihi infre- 16:10 He teareth me in his wrath, who hateth me: he

gnasheth upon me with his teeth; mine enemy sharpeneth his eyes upon me.

muit contra me dentibus suis hostis meus terribilibus oculis me intuitus est 16:11 aperuerunt super me ora sua exprob- 16:11 They have gaped upon me with their mouth; they

have smitten me upon the cheek reproachfully; they have gathered themselves together against me.

rantes percusserunt maxillam meam satiati sunt poenis meis 16:12 conclusit me Deus apud iniquum et manibus impio- 16:12 God hath delivered me to the ungodly, and turned

me over into the hands of the wicked.

rum me tradidit 16:13 ego ille quondam opulentus repente 16:13 I was at ease, but he hath broken me asunder: he

hath also taken me by my neck, and shaken me to pieces, and set me up for his mark.

contritus sum tenuit ceruicem meam confregit me et posuit sibi quasi in signum 16:14 circumdedit me lanceis suis con- 16:14 His archers compass me round about, he cleaveth

my reins asunder, and doth not spare; he poureth out my gall upon the ground.

uulnerauit lumbos meos non pepercit et effudit in terra uiscera mea 16:15 concidit me uulnere super uulnus inruit in 16:15 He breaketh me with breach upon breach, he run-

neth upon me like a giant.

me quasi gigans 16:16 saccum consui super cutem meam et 16:16 I have sewed sackcloth upon my skin, and defiled

my horn in the dust.

operui cinere cornu meum 16:17 facies mea intumuit a fletu 16:17 My face is foul with weeping, and on my eyelids is

the shadow of death;

et palpebrae meae caligauerunt 16:18 haec passus sum abs- 16:18 Not for any injustice in mine hands: also my prayer

is pure.

que iniquitate manus meae cum haberem mundas ad Deum preces 16:19 terra ne operias sanguinem meum neque in- 16:19 O earth, cover not thou my blood, and let my cry

have no place.

ueniat locum in te latendi clamor meus 16:20 ecce enim in 16:20 Also now, behold, my witness is in heaven, and my

record is on high.

caelo testis meus et conscius meus in excelsis 16:21 uer- 16:21 My friends scorn me: but mine eye poureth out tears

unto God.

bosi mei amici mei ad Deum stillat oculus meus 16:22 at- 16:22 O that one might plead for a man with God, as a

man pleadeth for his neighbour!

que utinam sic iudicaretur uir cum Deo quomodo iudicatur 16:23 When a few years are come, then I shall go the way filius hominis cum collega suo 16:23 ecce enim breues anni

whence I shall not return.

transeunt et semitam per quam non reuertar ambulo

17:1 My breath is corrupt, my days are extinct, the graves 17:1 spiritus meus adtenuabitur dies mei breuiabuntur et

are ready for me.

17:2 Are there not mockers with me? and doth not mine solum mihi superest sepulchrum 17:2 non peccaui et in

eye continue in their provocation?

17:3 Lay down now, put me in a surety with thee; who is amaritudinibus moratur oculus meus 17:3 libera me et pone

he that will strike hands with me? 17:4 For thou hast hid their heart from understanding: iuxta te et cuiusuis manus pugnet contra me 17:4 cor eotherefore shalt thou not exalt them.

rum longe fecisti a disciplina et propterea non exaltabun-

17:5 He that speaketh flattery to his friends, even the eyes tur 17:5 praedam pollicetur sociis et oculi filiorum eius de-

of his children shall fail.

17:6 He hath made me also a byword of the people; and ficient 17:6 posuit me quasi in prouerbium uulgi et exem-

aforetime I was as a tabret.

17:7 Mine eye also is dim by reason of sorrow, and all my plum sum coram eis 17:7 caligauit ab indignatione oculus

members are as a shadow. 17:8 Upright men shall be astonied at this, and the inno- meus et membra mea quasi in nihili redacta sunt 17:8 stucent shall stir up himself against the hypocrite.

pebunt iusti super hoc et innocens contra hypocritam susci-

17:9 The righteous also shall hold on his way, and he that tabitur 17:9 et tenebit iustus uiam suam et mundis manibus

hath clean hands shall be stronger and stronger.

17:10 But as for you all, do ye return, and come now: for addet fortitudinem 17:10 igitur uos omnes conuertimini et

I cannot find one wise man among you.

17:11 My days are past, my purposes are broken off, even uenite et non inueniam in uobis ullum sapientem 17:11 dies

the thoughts of my heart.

mei transierunt cogitationes meae dissipatae sunt torquentes

17:12 They change the night into day: the light is short cor meum 17:12 noctem uerterunt in diem et rursum post because of darkness.

17:13 If I wait, the grave is mine house: I have made my tenebras spero lucem 17:13 si sustinuero infernus domus

bed in the darkness. 17:14 I have said to corruption, Thou art my father: to the mea est in tenebris straui lectulum meum 17:14 putredini worm, Thou art my mother, and my sister.

17:15 And where is now my hope? as for my hope, who dixi pater meus es mater mea et soror mea uermibus 17:15

shall see it?

ubi est ergo nunc praestolatio mea et patientiam meam quis

17:16 They shall go down to the bars of the pit, when our considerat 17:16 in profundissimum infernum descendent

rest together is in the dust.

omnia mea putasne saltim ibi erit requies mihi

18:1 Then answered Bildad the Shuhite, and said, 18:1 respondens autem Baldad Suites dixit 18:2 usque ad

18:2 How long will it be ere ye make an end of words? mark, and afterwards we will speak.

quem finem uerba iactabitis intellegite prius et sic loqua-

18:3 Wherefore are we counted as beasts, and reputed vile mur 18:3 quare reputati sumus ut iumenta et sorduimus coin your sight? 18:4 He teareth himself in his anger: shall the earth be ram uobis 18:4 qui perdis animam tuam in furore tuo numforsaken for thee? and shall the rock be removed out of his place?

quid propter te derelinquetur terra et transferentur rupes de

18:5 Yea, the light of the wicked shall be put out, and the loco suo 18:5 nonne lux impii extinguetur nec splendebit

spark of his fire shall not shine.

18:6 The light shall be dark in his tabernacle, and his can- flamma ignis eius 18:6 lux obtenebrescet in tabernaculo il-

dle shall be put out with him.

18:7 The steps of his strength shall be straitened, and his lius et lucerna quae super eum est extinguetur 18:7 artabun-

own counsel shall cast him down.

tur gressus uirtutis eius et praecipitabit eum consilium suum

18:8 For he is cast into a net by his own feet, and he 18:8 inmisit enim in rete pedes suos et in maculis eius ambu-

walketh upon a snare.

lat 18:9 tenebitur planta illius laqueo et exardescet contra 18:9 The gin shall take him by the heel, and the robber

shall prevail against him.

eum sitis 18:10 abscondita est in terra pedica eius et de- 18:10 The snare is laid for him in the ground, and a trap

for him in the way.

cipula illius super semitam 18:11 undique terrebunt eum 18:11 Terrors shall make him afraid on every side, and

shall drive him to his feet.

formidines et inuoluent pedes eius 18:12 adtenuetur fame 18:12 His strength shall be hungerbitten, and destruction

shall be ready at his side.

robur eius et inedia inuadat costas illius 18:13 deuoret pul- 18:13 It shall devour the strength of his skin: even the

firstborn of death shall devour his strength.

chritudinem cutis eius consumat brachia illius primogenita mors 18:14 auellatur de tabernaculo suo fiducia eius et cal- 18:14 His confidence shall be rooted out of his tabernacle,

and it shall bring him to the king of terrors.

cet super eum quasi rex interitus 18:15 habitent in taberna- 18:15 It shall dwell in his tabernacle, because it is none of

his: brimstone shall be scattered upon his habitation.

culo illius socii eius qui non est aspergatur in tabernaculo eius sulphur 18:16 deorsum radices eius siccentur sursum 18:16 His roots shall be dried up beneath, and above shall

his branch be cut off.

autem adteratur messis eius 18:17 memoria illius pereat de 18:17 His remembrance shall perish from the earth, and

he shall have no name in the street.

terra et non celebretur nomen eius in plateis 18:18 expel- 18:18 He shall be driven from light into darkness, and

chased out of the world.

let eum de luce in tenebras et de orbe transferet eum 18:19 18:19 He shall neither have son nor nephew among his

people, nor any remaining in his dwellings.

non erit semen eius neque progenies in populo suo nec ullae reliquiae in regionibus eius 18:20 in die eius stupebunt 18:20 They that come after him shall be astonied at his

day, as they that went before were affrighted.

nouissimi et primos inuadet horror 18:21 haec sunt ergo ta- 18:21 Surely such are the dwellings of the wicked, and this

is the place of him that knoweth not God.

bernacula iniqui et iste locus eius qui ignorat Deum

19:1 respondens autem Iob dixit 19:2 usquequo adfligi- 19:1 Then Job answered and said,

19:2 How long will ye vex my soul, and break me in pieces with words?

tis animam meam et adteritis me sermonibus 19:3 en de-

19:3 These ten times have ye reproached me: ye are not ashamed that ye make yourselves strange to me.

cies confunditis me et non erubescitis opprimentes me 19:4

19:4 And be it indeed that I have erred, mine error remaineth with myself.

nempe et si ignoraui mecum erit ignorantia mea 19:5 at uos _{19:5 If indeed ye will magnify yourselves against me, and}

plead against me my reproach:

contra me erigimini et arguitis me obprobriis meis 19:6 sal- 19:6 Know now that God hath overthrown me, and hath

compassed me with his net.

tim nunc intellegite quia Deus non aequo iudicio adflixerit me et flagellis suis me cinxerit 19:7 ecce clamabo uim pa- 19:7 Behold, I cry out of wrong, but I am not heard: I cry

aloud, but there is no judgment.

tiens et nemo audiet uociferabor et non est qui iudicet 19:8 19:8 He hath fenced up my way that I cannot pass, and he

hath set darkness in my paths.

semitam meam circumsepsit et transire non possum et in calle meo tenebras posuit 19:9 spoliauit me gloria mea et 19:9 He hath stripped me of my glory, and taken the crown

from my head.

abstulit coronam de capite meo 19:10 destruxit me undique 19:10 He hath destroyed me on every side, and I am gone:

and mine hope hath he removed like a tree.

et pereo et quasi euulsae arbori abstulit spem meam 19:11 19:11 He hath also kindled his wrath against me, and he

counteth me unto him as one of his enemies.

iratus est contra me furor eius et sic me habuit quasi hostem suum 19:12 simul uenerunt latrones eius et fecerunt 19:12 His troops come together, and raise up their way

against me, and encamp round about my tabernacle.

sibi uiam per me et obsederunt in gyro tabernaculum meum 19:13 fratres meos longe fecit a me et noti mei quasi alieni 19:13 He hath put my brethren far from me, and mine ac-

quaintance are verily estranged from me.

19:14 My kinsfolk have failed, and my familiar friends recesserunt a me 19:14 dereliquerunt me propinqui mei et

have forgotten me.

19:15 They that dwell in mine house, and my maids, count qui me nouerant obliti sunt mei 19:15 inquilini domus meae

me for a stranger: I am an alien in their sight.

et ancillae meae sicut alienum habuerunt me et quasi per-

19:16 I called my servant, and he gave me no answer; I egrinus fui in oculis eorum 19:16 seruum meum uocaui et

intreated him with my mouth.

19:17 My breath is strange to my wife, though I intreated non respondit ore proprio deprecabar illum 19:17 halitum

for the children’s sake of mine own body. 19:18 Yea, young children despised me; I arose, and they meum exhorruit uxor mea et orabam filios uteri mei 19:18 spake against me.

stulti quoque despiciebant me et cum ab eis recessissem de-

19:19 All my inward friends abhorred me: and they whom trahebant mihi 19:19 abominati sunt me quondam consilia-

I loved are turned against me.

19:20 My bone cleaveth to my skin and to my flesh, and I rii mei et quem maxime diligebam auersatus est me 19:20

am escaped with the skin of my teeth.

pelli meae consumptis carnibus adhesit os meum et derelicta

19:21 Have pity upon me, have pity upon me, O ye my sunt tantummodo labia circa dentes meos 19:21 miseremini

friends; for the hand of God hath touched me.

mei miseremini mei saltim uos amici mei quia manus Do-

19:22 Why do ye persecute me as God, and are not satis- mini tetigit me 19:22 quare persequimini me sicut Deus et

fied with my flesh?

19:23 Oh that my words were now written! oh that they carnibus meis saturamini 19:23 quis mihi tribuat ut scriban-

were printed in a book!

19:24 That they were graven with an iron pen and lead in tur sermones mei quis mihi det ut exarentur in libro 19:24

the rock for ever!

stilo ferreo et plumbi lammina uel certe sculpantur in silice

19:25 For I know that my redeemer liveth, and that he shall 19:25 scio enim quod redeeptor meus uiuat et in nouissimo stand at the latter day upon the earth:

19:26 And though after my skin worms destroy this body, de terra surrecturus sim 19:26 et rursum circumdabor pelle

yet in my flesh shall I see God: 19:27 Whom I shall see for myself, and mine eyes shall mea et in carne mea uidebo Deum 19:27 quem uisurus sum behold, and not another; though my reins be consumed within me.

ego ipse et oculi mei conspecturi sunt et non alius reposita

19:28 But ye should say, Why persecute we him, seeing est haec spes mea in sinu meo 19:28 quare ergo nunc dithe root of the matter is found in me?

citis persequamur eum et radicem uerbi inueniamus contra

19:29 Be ye afraid of the sword: for wrath bringeth the eum 19:29 fugite ergo a facie gladii quoniam ultor iniquita-

punishments of the sword, that ye may know there is a judgment.

tum gladius est et scitote esse iudicium

20:1 Then answered Zophar the Naamathite, and said, 20:1 respondens autem Sophar Naamathites dixit 20:2 id-

20:2 Therefore do my thoughts cause me to answer, and for this I make haste.

circo cogitationes meae uariae succedunt sibi et mens in di-

20:3 I have heard the check of my reproach, and the spirit uersa rapitur 20:3 doctrinam qua me arguis audiam et spi-

of my understanding causeth me to answer.

20:4 Knowest thou not this of old, since man was placed ritus intellegentiae meae respondebit mihi 20:4 hoc scio a

upon earth,

20:5 That the triumphing of the wicked is short, and the principio ex quo positus est homo super terram 20:5 quod

joy of the hypocrite but for a moment?

laus impiorum breuis sit et gaudium hypocritae ad instar

20:6 Though his excellency mount up to the heavens, and puncti 20:6 si ascenderit usque ad caelum superbia eius et

his head reach unto the clouds;

20:7 Yet he shall perish for ever like his own dung: they caput eius nubes tetigerit 20:7 quasi sterquilinium in fine

which have seen him shall say, Where is he?

20:8 He shall fly away as a dream, and shall not be found: perdetur et qui eum uiderant dicent ubi est 20:8 uelut som-

yea, he shall be chased away as a vision of the night.

nium auolans non inuenietur transiet sicut uisio nocturna 20:9 oculus qui eum uiderat non uidebit neque ultra intue- 20:9 The eye also which saw him shall see him no more;

neither shall his place any more behold him.

bitur eum locus suus 20:10 filii eius adterentur egestate et 20:10 His children shall seek to please the poor, and his

hands shall restore their goods.

manus illius reddent ei dolorem suum 20:11 ossa eius im- 20:11 His bones are full of the sin of his youth, which shall

lie down with him in the dust.

plebuntur uitiis adulescentiae eius et cum eo in puluerem dormient 20:12 cum enim dulce fuerit in ore eius malum 20:12 Though wickedness be sweet in his mouth, though

he hide it under his tongue;

abscondet illud sub lingua sua 20:13 parcet illi et non de- 20:13 Though he spare it, and forsake it not; but keep it

still within his mouth:

relinquet illud et celabit in gutture suo 20:14 panis eius in 20:14 Yet his meat in his bowels is turned, it is the gall of

asps within him.

utero illius uertetur in fel aspidum intrinsecus 20:15 diuitias 20:15 He hath swallowed down riches, and he shall vomit

them up again: God shall cast them out of his belly.

quas deuorauit euomet et de uentre illius extrahet eas Deus 20:16 caput aspidum suget occidet eum lingua uiperae 20:17 20:16 He shall suck the poison of asps: the viper’s tongue

shall slay him.

20:17 He shall not see the rivers, the floods, the brooks of honey and butter.

non uideat riuulos fluminis torrentes mellis et butyri 20:18

20:18 That which he laboured for shall he restore, and shall not swallow it down: according to his substance shall the restitution be, and he shall not rejoice therein.

luet quae fecit omnia nec tamen consumetur iuxta multitudinem adinuentionum suarum sic et sustinebit 20:19 quoniam 20:19 Because he hath oppressed and hath forsaken the

poor; because he hath violently taken away an house which he builded not;

confringens nudauit pauperes domum rapuit et non aedificauit eam 20:20 nec est satiatus uenter eius et cum habuerit 20:20 Surely he shall not feel quietness in his belly, he

shall not save of that which he desired.

quae cupierat possidere non poterit 20:21 non remansit de 20:21 There shall none of his meat be left; therefore shall

no man look for his goods.

cibo eius et propterea nihil permanebit de bonis eius 20:22 20:22 In the fulness of his sufficiency he shall be in straits:

every hand of the wicked shall come upon him.

cum satiatus fuerit artabitur aestuabit et omnis dolor inruet in eum 20:23 utinam impleatur uenter eius ut emittat in eum 20:23 When he is about to fill his belly, God shall cast

the fury of his wrath upon him, and shall rain it upon him while he is eating.

iram furoris sui et pluat super illum bellum suum 20:24 fu-

20:24 He shall flee from the iron weapon, and the bow of steel shall strike him through.

giet arma ferrea et inruet in arcum aereum 20:25 eductus _{20:25 It is drawn, and cometh out of the body; yea, the}

glittering sword cometh out of his gall: terrors are upon him.

et egrediens de uagina sua et fulgurans in amaritudine sua uadent et uenient super eum horribiles 20:26 omnes teneb- 20:26 All darkness shall be hid in his secret places: a fire

not blown shall consume him; it shall go ill with him that is left in his tabernacle.

rae absconditae sunt in occultis eius deuorabit eum ignis qui non succenditur adfligetur relictus in tabernaculo suo 20:27 20:27 The heaven shall reveal his iniquity; and the earth

shall rise up against him.

reuelabunt caeli iniquitatem eius et terra consurget aduersus

eum 20:28 apertum erit germen domus illius detrahetur in 20:28 The increase of his house shall depart, and his goods

shall flow away in the day of his wrath.

die furoris Dei 20:29 haec est pars hominis impii a Deo et 20:29 This is the portion of a wicked man from God, and

the heritage appointed unto him by God.

hereditas uerborum eius a Domino

21:1 respondens autem Iob dixit 21:2 audite quaeso ser- 21:1 But Job answered and said,

21:2 Hear diligently my speech, and let this be your consolations.

mones meos et agetis paenitentiam 21:3 sustinete me ut et

21:3 Suffer me that I may speak; and after that I have spoken, mock on.

ego loquar et post mea si uidebitur uerba ridete 21:4 num-

21:4 As for me, is my complaint to man? and if it were so, why should not my spirit be troubled?

quid contra hominem disputatio mea est ut merito non de-

21:5 Mark me, and be astonished, and lay your hand upon beam contristari 21:5 adtendite me et obstupescite et su-

your mouth.

21:6 Even when I remember I am afraid, and trembling perponite digitum ori uestro 21:6 et ego quando recorda-

taketh hold on my flesh.

21:7 Wherefore do the wicked live, become old, yea, are tus fuero pertimesco et concutit carnem meam tremor 21:7

mighty in power?

quare ergo impii uiuunt subleuati sunt confortatique diuitiis

21:8 Their seed is established in their sight with them, and 21:8 semen eorum permanet coram eis propinquorum turba their offspring before their eyes. 21:9 Their houses are safe from fear, neither is the rod of et nepotum in conspectu eorum 21:9 domus eorum securae God upon them. 21:10 Their bull gendereth, and faileth not; their cow cal- sunt et pacatae et non est uirga Dei super illos 21:10 bos veth, and casteth not her calf.

eorum concepit et non abortit uacca peperit et non est pri-

21:11 They send forth their little ones like a flock, and uata fetu suo 21:11 egrediuntur quasi greges paruuli eorum

their children dance.

21:12 They take the timbrel and harp, and rejoice at the et infantes eorum exultant lusibus 21:12 tenent tympanum

sound of the organ.

21:13 They spend their days in wealth, and in a moment et citharam et gaudent ad sonitum organi 21:13 ducunt in

go down to the grave.

21:14 Therefore they say unto God, Depart from us; for bonis dies suos et in puncto ad inferna descendunt 21:14

we desire not the knowledge of thy ways.

qui dixerunt Deo recede a nobis et scientiam uiarum tua-

21:15 What is the Almighty, that we should serve him? rum nolumus 21:15 quid est Omnipotens ut seruiamus ei

and what profit should we have, if we pray unto him?

21:16 Lo, their good is not in their hand: the counsel of et quid nobis prodeet si orauerimus illum 21:16 uerumta-

the wicked is far from me.

men quia non sunt in manu eorum bona sua consilium im-

21:17 How oft is the candle of the wicked put out! and piorum longe sit a me 21:17 quotiens lucerna impiorum ex-

how oft cometh their destruction upon them! God distributeth sorrows in his anger.

tinguetur et superueniet eis inundatio et dolores diuidet fu-

21:18 They are as stubble before the wind, and as chaff roris sui 21:18 erunt sicut paleae ante faciem uenti et sicut

that the storm carrieth away.

21:19 God layeth up his iniquity for his children: he re- fauilla quam turbo dispergit 21:19 Deus seruabit filiis illius

wardeth him, and he shall know it. 21:20 His eyes shall see his destruction, and he shall drink dolorem patris et cum reddiderit tunc sciet 21:20 uidebunt of the wrath of the Almighty.

oculi eius interfectionem suam et de furore Omnipotentis bi-

21:21 For what pleasure hath he in his house after him, bet 21:21 quid enim ad eum pertinet de domo sua post se et

when the number of his months is cut off in the midst? 21:22 Shall any teach God knowledge? seeing he judgeth si numerus mensuum eius dimidietur 21:22 numquid Dethose that are high.

21:23 One dieth in his full strength, being wholly at ease um quispiam docebit scientiam qui excelsos iudicat 21:23

and quiet.

21:24 His breasts are full of milk, and his bones are moist- iste moritur robustus et sanus diues et felix 21:24 uiscera

ened with marrow.

21:25 And another dieth in the bitterness of his soul, and eius plena sunt adipe et medullis ossa illius inrigantur 21:25

never eateth with pleasure.

alius uero moritur in amaritudine animae absque ullis opibus

21:26 They shall lie down alike in the dust, and the worms 21:26 et tamen simul in puluerem dormient et uermes opeshall cover them.

21:27 Behold, I know your thoughts, and the devices rient eos 21:27 certe noui cogitationes uestras et sententias

which ye wrongfully imagine against me.

21:28 For ye say, Where is the house of the prince? and contra me iniquas 21:28 dicitis enim ubi est domus princi-

where are the dwelling places of the wicked?

21:29 Have ye not asked them that go by the way? and do pis et ubi tabernacula impiorum 21:29 interrogate quemli-

ye not know their tokens,

bet de uiatoribus et haec eadem eum intellegere cognoscetis 21:30 quia in diem perditionis seruabitur malus et ad diem fu- 21:30 That the wicked is reserved to the day of destruc-

tion? they shall be brought forth to the day of wrath.

roris ducitur 21:31 quis arguet coram eo uiam eius et quae 21:31 Who shall declare his way to his face? and who

shall repay him what he hath done?

fecit quis reddet illi 21:32 ipse ad sepulchra ducetur et in 21:32 Yet shall he be brought to the grave, and shall re-

main in the tomb.

congerie mortuorum uigilabit 21:33 dulcis fuit glareis Co- 21:33 The clods of the valley shall be sweet unto him, and

every man shall draw after him, as there are innumerable before him.

cyti et post se omnem hominem trahet et ante se innumerabiles 21:34 quomodo igitur consolamini me frustra cum 21:34 How then comfort ye me in vain, seeing in your

answers there remaineth falsehood?

responsio uestra repugnare ostensa sit ueritati

22:1 respondens autem Eliphaz Themanites dixit 22:2 22:1 Then Eliphaz the Temanite answered and said,

22:2 Can a man be profitable unto God, as he that is wise may be profitable unto himself?

numquid Deo conparari potest homo etiam cum perfectae fuerit scientiae 22:3 quid prodeet Deo si iustus fueris aut 22:3 Is it any pleasure to the Almighty, that thou art righ-

teous? or is it gain to him, that thou makest thy ways perfect?

quid ei confers si inmaculata fuerit uia tua 22:4 numquid

22:4 Will he reprove thee for fear of thee? will he enter with thee into judgment?

timens arguet te et ueniet tecum in iudicium 22:5 et non 22:5 Is not thy wickedness great? and thine iniquities infi-

nite?

propter malitiam tuam plurimam et infinitas iniquitates tuas 22:6 abstulisti enim pignus fratrum tuorum sine causa et nu- 22:6 For thou hast taken a pledge from thy brother for

nought, and stripped the naked of their clothing.

dos spoliasti uestibus 22:7 aquam lasso non dedisti et esu- 22:7 Thou hast not given water to the weary to drink, and

thou hast withholden bread from the hungry.

rienti subtraxisti panem 22:8 in fortitudine brachii tui pos- 22:8 But as for the mighty man, he had the earth; and the

honourable man dwelt in it.

sidebas terram et potentissimus obtinebas eam 22:9 uiduas 22:9 Thou hast sent widows away empty, and the arms of

the fatherless have been broken.

dimisisti uacuas et lacertos pupillorum comminuisti 22:10 22:10 Therefore snares are round about thee, and sudden

fear troubleth thee;

propterea circumdatus es laqueis et conturbat te formido su-

bita 22:11 et putabas te tenebras non uisurum et impetu 22:11 Or darkness, that thou canst not see; and abundance

of waters cover thee.

aquarum inundantium non oppressurum 22:12 an cogitas 22:12 Is not God in the height of heaven? and behold the

height of the stars, how high they are!

quod Deus excelsior caelo et super stellarum uertices sublimetur 22:13 et dicis quid enim nouit Deus et quasi per cali- 22:13 And thou sayest, How doth God know? can he judge

through the dark cloud?

ginem iudicat 22:14 nubes latibulum eius nec nostra consi- 22:14 Thick clouds are a covering to him, that he seeth

not; and he walketh in the circuit of heaven.

derat et circa cardines caeli perambulat 22:15 numquid se- 22:15 Hast thou marked the old way which wicked men

have trodden?

mitam saeculorum custodire cupis quam calcauerunt uiri iniqui 22:16 qui sublati sunt ante tempus suum et fluuius sub- 22:16 Which were cut down out of time, whose foundation

was overflown with a flood:

uertit fundamentum eorum 22:17 qui dicebant Deo recede 22:17 Which said unto God, Depart from us: and what can

the Almighty do for them?

a nobis et quasi nihil possit facere Omnipotens aestimabant

eum 22:18 cum ille implesset domos eorum bonis quorum 22:18 Yet he filled their houses with good things: but the

counsel of the wicked is far from me.

sententia procul sit a me 22:19 uidebunt iusti et laetabuntur 22:19 The righteous see it, and are glad: and the innocent

laugh them to scorn.

et innocens subsannabit eos 22:20 nonne succisa est erectio 22:20 Whereas our substance is not cut down, but the rem-

nant of them the fire consumeth.

22:21 Acquaint now thyself with him, and be at peace: eorum et reliquias eorum deuorauit ignis 22:21 adquiesce

thereby good shall come unto thee.

igitur ei et habeto pacem et per haec habebis fructus optimos

22:22 Receive, I pray thee, the law from his mouth, and 22:22 suscipe ex ore illius legem et pone sermones eius in

lay up his words in thine heart. 22:23 If thou return to the Almighty, thou shalt be built up, corde tuo 22:23 si reuersus fueris ad Omnipotentem aedifithou shalt put away iniquity far from thy tabernacles.

22:24 Then shalt thou lay up gold as dust, and the gold of caberis et longe facies iniquitatem a tabernaculo tuo 22:24

Ophir as the stones of the brooks.

22:25 Yea, the Almighty shall be thy defence, and thou dabit pro terra silicem et pro silice torrentes aureos 22:25

shalt have plenty of silver.

eritque Omnipotens contra hostes tuos et argentum coacer-

22:26 For then shalt thou have thy delight in the Almighty, uabitur tibi 22:26 tunc super Omnipotentem deliciis afflues

and shalt lift up thy face unto God.

22:27 Thou shalt make thy prayer unto him, and he shall et eleuabis ad Deum faciem tuam 22:27 rogabis eum et ex-

hear thee, and thou shalt pay thy vows.

22:28 Thou shalt also decree a thing, and it shall be estab- audiet te et uota tua reddes 22:28 decernes rem et ueniet tibi

lished unto thee: and the light shall shine upon thy ways.

22:29 When men are cast down, then thou shalt say, There et in uiis tuis splendebit lumen 22:29 qui enim humiliatus

is lifting up; and he shall save the humble person.

fuerit erit in gloria et qui inclinauerit oculos suos ipse salua-

22:30 He shall deliver the island of the innocent: and it is bitur 22:30 saluabitur innocens saluabitur autem munditia

delivered by the pureness of thine hands.

manuum suarum

23:1 Then Job answered and said, 23:1 respondens autem Iob dixit 23:2 nunc quoque in

23:2 Even to day is my complaint bitter: my stroke is heavier than my groaning.

amaritudine est sermo meus et manus plagae meae adgra-

23:3 Oh that I knew where I might find him! that I might uata est super gemitum meum 23:3 quis mihi tribuat ut co-

come even to his seat!

gnoscam et inueniam illum et ueniam usque ad solium eius

23:4 I would order my cause before him, and fill my mouth 23:4 ponam coram eo iudicium et os meum replebo increpa-

with arguments.

23:5 I would know the words which he would answer me, tionibus 23:5 ut sciam uerba quae mihi respondeat et intel-

and understand what he would say unto me.

23:6 Will he plead against me with his great power? No; legam quid loquatur mihi 23:6 nolo multa fortitudine con-

but he would put strength in me.

23:7 There the righteous might dispute with him; so tendat mecum nec magnitudinis suae mole me premat 23:7

should I be delivered for ever from my judge.

proponat aequitatem contra me et perueniat ad uictoriam iu-

23:8 Behold, I go forward, but he is not there; and back- dicium meum 23:8 si ad orientem iero non apparet si ad ward, but I cannot perceive him:

23:9 On the left hand, where he doth work, but I cannot occidentem non intellegam eum 23:9 si ad sinistram quid

behold him: he hideth himself on the right hand, that I cannot see him:

agat non adprehendam eum si me uertam ad dextram non ui-

23:10 But he knoweth the way that I take: when he hath debo illum 23:10 ipse uero scit uiam meam et probauit me

tried me, I shall come forth as gold.

23:11 My foot hath held his steps, his way have I kept, and quasi aurum quod per ignem transit 23:11 uestigia eius se-

not declined.

cutus est pes meus uiam eius custodiui et non declinaui ex ea

23:12 Neither have I gone back from the commandment 23:12 a mandatis labiorum eius non recessi et in sinu meo

of his lips; I have esteemed the words of his mouth more than my necessary food.

abscondi uerba oris eius 23:13 ipse enim solus est et nemo

23:13 But he is in one mind, and who can turn him? and what his soul desireth, even that he doeth.

auertere potest cogitationem eius et anima eius quodcumque

23:14 For he performeth the thing that is appointed for me: uoluerit hoc facit 23:14 cum expleuerit in me uoluntatem and many such things are with him.

suam et alia multa similia praesto sunt ei 23:15 et idcirco a 23:15 Therefore am I troubled at his presence: when I con-

sider, I am afraid of him.

facie eius turbatus sum et considerans eum timore sollicitor 23:16 Deus molliuit cor meum et Omnipotens conturbauit me 23:16 For God maketh my heart soft, and the Almighty

troubleth me:

23:17 non enim perii propter inminentes tenebras nec faciem 23:17 Because I was not cut off before the darkness, nei-

ther hath he covered the darkness from my face.

meam operuit caligo

24:1 ab Omnipotente non sunt abscondita tempora qui au- 24:1 Why, seeing times are not hidden from the Almighty,

do they that know him not see his days?

tem nouerunt eum ignorant dies illius 24:2 alii terminos 24:2 Some remove the landmarks; they violently take

away flocks, and feed thereof.

transtulerunt diripuerunt greges et pauerunt eos 24:3 asi- 24:3 They drive away the ass of the fatherless, they take

the widow’s ox for a pledge.

num pupillorum abigerunt et abstulerunt pro pignore bouem uiduae 24:4 subuerterunt pauperum uiam et oppres- 24:4 They turn the needy out of the way: the poor of the

earth hide themselves together.

serunt pariter mansuetos terrae 24:5 alii quasi onagri in de- 24:5 Behold, as wild asses in the desert, go they forth to

their work; rising betimes for a prey: the wilderness yieldeth food for them and for their children.

serto egrediuntur ad opus suum uigilantesque ad praedam praeparant panem liberis 24:6 agrum non suum demetunt 24:6 They reap every one his corn in the field: and they

gather the vintage of the wicked.

et uineam eius quem ui oppresserunt uindemiant 24:7 nu- 24:7 They cause the naked to lodge without clothing, that

they have no covering in the cold.

dos dimittunt homines indumenta tollentes quibus non est operimentum in frigore 24:8 quos imbres montium rigant et 24:8 They are wet with the showers of the mountains, and

embrace the rock for want of a shelter.

non habentes uelamen amplexantur lapides 24:9 uim fece- 24:9 They pluck the fatherless from the breast, and take a

pledge of the poor.

runt depraedantes pupillos et uulgum pauperem spoliauerunt 24:10 nudis et incedentibus absque uestitu et esurientibus tu- 24:10 They cause him to go naked without clothing, and

they take away the sheaf from the hungry;

lerunt spicas 24:11 inter aceruos eorum meridiati sunt qui 24:11 Which make oil within their walls, and tread their

winepresses, and suffer thirst.

calcatis torcularibus sitiunt 24:12 de ciuitatibus fecerunt ui- 24:12 Men groan from out of the city, and the soul of the

wounded crieth out: yet God layeth not folly to them.

ros gemere et anima uulneratorum clamauit et Deus inultum abire non patitur 24:13 ipsi fuerunt rebelles luminis nescie- 24:13 They are of those that rebel against the light; they

know not the ways thereof, nor abide in the paths thereof.

runt uias eius nec reuersi sunt per semitas illius 24:14 mane 24:14 The murderer rising with the light killeth the poor

and needy, and in the night is as a thief.

primo consurgit homicida interficit egenum et pauperem per

noctem uero erit quasi fur 24:15 oculus adulteri obseruat 24:15 The eye also of the adulterer waiteth for the twilight,

saying, No eye shall see me: and disguiseth his face.

caliginem dicens non me uidebit oculus et operiet uultum suum 24:16 perfodit in tenebris domos sicut in die condixe- 24:16 In the dark they dig through houses, which they had

marked for themselves in the daytime: they know not the light.

rant sibi et ignorauerunt lucem 24:17 si subito apparuerit

24:17 For the morning is to them even as the shadow of death: if one know them, they are in the terrors of the shadow of death.

aurora arbitrantur umbram mortis et sic in tenebris quasi in luce ambulant 24:18 leuis est super faciem aquae maledicta 24:18 He is swift as the waters; their portion is cursed in

the earth: he beholdeth not the way of the vineyards.

sit pars eius in terra nec ambulet per uiam uinearum 24:19 24:19 Drought and heat consume the snow waters: so doth

the grave those which have sinned.

ad nimium calorem transeat ab aquis niuium et usque ad in-24:20 The womb shall forget him; the worm shall feed feros peccatum illius 24:20 obliuiscatur eius misericordia

sweetly on him; he shall be no more remembered; and wickedness shall be broken as a tree.

dulcedo illius uermes non sit in recordatione sed contera-

24:21 He evil entreateth the barren that beareth not: and tur quasi lignum infructuosum 24:21 pauit enim sterilem et

doeth not good to the widow.

24:22 He draweth also the mighty with his power: he quae non parit et uiduae bene non fecit 24:22 detraxit for-

riseth up, and no man is sure of life.

tes in fortitudine sua et cum steterit non credet uitae suae

24:23 Though it be given him to be in safety, whereon he 24:23 dedit ei Deus locum paenitentiae et ille abutitur eo in

resteth; yet his eyes are upon their ways.

24:24 They are exalted for a little while, but are gone and superbiam oculi autem eius sunt in uiis illius 24:24 eleuati

brought low; they are taken out of the way as all other, and cut off as the tops of the ears of corn.

sunt ad modicum et non subsistent et humiliabuntur sicut omnia et auferentur et sicut summitates spicarum conteren-

24:25 And if it be not so now, who will make me a liar, tur 24:25 quod si non est ita quis me potest arguere esse

and make my speech nothing worth?

mentitum et ponere ante Deum uerba mea

25:1 Then answered Bildad the Shuhite, and said, 25:1 respondens autem Baldad Suites dixit 25:2 potes-

25:2 Dominion and fear are with him, he maketh peace in his high places.

tas et terror apud eum est qui facit concordiam in sublimi-

25:3 Is there any number of his armies? and upon whom bus suis 25:3 numquid est numerus militum eius et super

doth not his light arise?

25:4 How then can man be justified with God? or how can quem non surget lumen illius 25:4 numquid iustificari po-

he be clean that is born of a woman?

test homo conparatus Deo aut apparere mundus natus de mu-

25:5 Behold even to the moon, and it shineth not; yea, the liere 25:5 ecce etiam luna non splendet et stellae non sunt

stars are not pure in his sight.

25:6 How much less man, that is a worm? and the son of mundae in conspectu eius 25:6 quanto magis homo putredo

man, which is a worm?

et filius hominis uermis

26:1 But Job answered and said, 26:1 respondens autem Iob dixit 26:2 cuius adiutor es

26:2 How hast thou helped him that is without power? how savest thou the arm that hath no strength?

numquid inbecilli et sustentas brachium eius qui non est

26:3 How hast thou counselled him that hath no wisdom? fortis 26:3 cui dedisti consilium forsitan illi qui non habet

and how hast thou plentifully declared the thing as it is?

26:4 To whom hast thou uttered words? and whose spirit sapientiam et prudentiam tuam ostendisti plurimam 26:4

came from thee?

quem docere uoluisti nonne eum qui fecit spiramen tuum

26:5 Dead things are formed from under the waters, and 26:5 ecce gigantes gemunt sub aquis et qui habitant cum eis

the inhabitants thereof.

26:6 Hell is naked before him, and destruction hath no 26:6 nudus est inferus coram illo et nullum est operimentum

covering.

26:7 He stretcheth out the north over the empty place, and perditioni 26:7 qui extendit aquilonem super uacuum et ad-

hangeth the earth upon nothing.

26:8 He bindeth up the waters in his thick clouds; and the pendit terram super nihili 26:8 qui ligat aquas in nubibus

cloud is not rent under them.

26:9 He holdeth back the face of his throne, and spreadeth suis ut non erumpant pariter deorsum 26:9 qui tenet uultum

his cloud upon it.

26:10 He hath compassed the waters with bounds, until solii sui et expandit super illud nebulam suam 26:10 termi-

the day and night come to an end.

num circumdedit aquis usque dum finiantur lux et tenebrae

26:11 The pillars of heaven tremble and are astonished at 26:11 columnae caeli contremescunt et pauent ad nutum eius

his reproof.

26:12 He divideth the sea with his power, and by his un- 26:12 in fortitudine illius repente maria congregata sunt et

derstanding he smiteth through the proud.

prudentia eius percussit superbum 26:13 spiritus eius orna- 26:13 By his spirit he hath garnished the heavens; his hand

hath formed the crooked serpent.

uit caelos et obsetricante manu eius eductus est coluber tor-

tuosus 26:14 ecce haec ex parte dicta sunt uiarum eius et 26:14 Lo, these are parts of his ways: but how little a portion is heard of him? but the thunder of his power who can understand?

cum uix paruam stillam sermonis eius audierimus quis poterit tonitruum magnitudinis illius intueri

27:1 addidit quoque Iob adsumens parabolam suam et dixit 27:1 Moreover Job continued his parable, and said, 27:2 uiuit Deus qui abstulit iudicium meum et Omnipotens 27:2 As God liveth, who hath taken away my judgment;

and the Almighty, who hath vexed my soul;

qui ad amaritudinem adduxit animam meam 27:3 quia do- 27:3 All the while my breath is in me, and the spirit of

God is in my nostrils;

nec superest halitus in me et spiritus Dei in naribus meis 27:4 non loquentur labia mea iniquitatem nec lingua mea me- 27:4 My lips shall not speak wickedness, nor my tongue

utter deceit.

ditabitur mendacium 27:5 absit a me ut iustos uos esse iu- 27:5 God forbid that I should justify you: till I die I will

not remove mine integrity from me.

dicem donec deficiam non recedam ab innocentia mea 27:6 27:6 My righteousness I hold fast, and will not let it go:

my heart shall not reproach me so long as I live.

iustificationem meam quam coepi tenere non deseram nec

enim reprehendit me cor meum in omni uita mea 27:7 sit 27:7 Let mine enemy be as the wicked, and he that riseth

up against me as the unrighteous.

ut impius inimicus meus et aduersarius meus quasi iniquus 27:8 quae enim spes est hypocritae si auare rapiat et non li- 27:8 For what is the hope of the hypocrite, though he hath

gained, when God taketh away his soul?

beret Deus animam eius 27:9 numquid clamorem eius Deus 27:9 Will God hear his cry when trouble cometh upon

him?

audiet cum uenerit super illum angustia 27:10 aut poterit in 27:10 Will he delight himself in the Almighty? will he

always call upon God?

Omnipotente delectari et inuocare Deum in omni tempore 27:11 docebo uos per manum Dei quae Omnipotens habeat 27:11 I will teach you by the hand of God: that which is

with the Almighty will I not conceal.

nec abscondam 27:12 ecce uos omnes nostis et quid sine 27:12 Behold, all ye yourselves have seen it; why then are

ye thus altogether vain?

causa uana loquimini 27:13 haec est pars hominis impii 27:13 This is the portion of a wicked man with God, and

the heritage of oppressors, which they shall receive of the Almighty.

apud Deum et hereditas uiolentorum quam ab Omnipotente suscipient 27:14 si multiplicati fuerint filii eius in gladio 27:14 If his children be multiplied, it is for the sword: and

his offspring shall not be satisfied with bread.

erunt et nepotes eius non saturabuntur pane 27:15 qui re- 27:15 Those that remain of him shall be buried in death:

and his widows shall not weep.

liqui fuerint ex eo sepelientur in interitu et uiduae illius non plorabunt 27:16 si conportauerit quasi terram argentum et 27:16 Though he heap up silver as the dust, and prepare

raiment as the clay;

sicut lutum praeparauerit uestimenta 27:17 praeparabit qui- 27:17 He may prepare it, but the just shall put it on, and

the innocent shall divide the silver.

dem sed iustus uestietur illis et argentum innocens diuidet 27:18 aedificauit sicut tinea domum suam et sicut custos fecit 27:18 He buildeth his house as a moth, and as a booth that

the keeper maketh.

umbraculum 27:19 diues cum dormierit nihil secum auferet 27:19 The rich man shall lie down, but he shall not be

gathered: he openeth his eyes, and he is not.

aperit oculos suos et nihil inueniet 27:20 adprehendit eum 27:20 Terrors take hold on him as waters, a tempest

stealeth him away in the night.

quasi aqua inopia nocte opprimet eum tempestas 27:21 tol- 27:21 The east wind carrieth him away, and he departeth:

and as a storm hurleth him out of his place.

let eum uentus urens et auferet et uelut turbo rapiet eum de

27:22 For God shall cast upon him, and not spare: he loco suo 27:22 et mittet super eum et non parcet de manu

would fain flee out of his hand.

27:23 Men shall clap their hands at him, and shall hiss him eius fugiens fugiet 27:23 stringet super eum manus suas et

out of his place.

sibilabit super illum intuens locum eius

28:1 Surely there is a vein for the silver, and a place for 28:1 habet argentum uenarum suarum principia et auro lo-

gold where they fine it.

28:2 Iron is taken out of the earth, and brass is molten out cus est in quo conflatur 28:2 ferrum de terra tollitur et lapis

of the stone.

28:3 He setteth an end to darkness, and searcheth out all solutus calore in aes uertitur 28:3 tempus posuit tenebris et

perfection: the stones of darkness, and the shadow of death.

uniuersorum finem ipse considerat lapidem quoque caligi-

28:4 The flood breaketh out from the inhabitant; even the nis et umbram mortis 28:4 diuidit torrens a populo pereg-

waters forgotten of the foot: they are dried up, they are gone away from men.

rinante eos quos oblitus est pes egentis hominum et inuios

28:5 As for the earth, out of it cometh bread: and under it 28:5 terra de qua oriebatur panis in loco suo igne subuersa

is turned up as it were fire.

28:6 The stones of it are the place of sapphires: and it hath est 28:6 locus sapphyri lapides eius et glebae illius aurum

dust of gold.

28:7 There is a path which no fowl knoweth, and which 28:7 semitam ignorauit auis nec intuitus est oculus uulturis

the vulture’s eye hath not seen:

28:8 The lion’s whelps have not trodden it, nor the fierce 28:8 non calcauerunt eam filii institorum nec pertransiuit per

lion passed by it.

28:9 He putteth forth his hand upon the rock; he overtur- eam leaena 28:9 ad silicem extendit manum suam subuertit

neth the mountains by the roots.

28:10 He cutteth out rivers among the rocks; and his eye a radicibus montes 28:10 in petris riuos excidit et omne pre-

seeth every precious thing.

28:11 He bindeth the floods from overflowing; and the tiosum uidit oculus eius 28:11 profunda quoque fluuiorum

thing that is hid bringeth he forth to light.

28:12 But where shall wisdom be found? and where is the scrutatus est et abscondita produxit in lucem 28:12 sapien-

place of understanding?

28:13 Man knoweth not the price thereof; neither is it tia uero ubi inuenitur et quis est locus intellegentiae 28:13

found in the land of the living.

nescit homo pretium eius nec inuenitur in terra suauiter ui-

28:14 The depth saith, It is not in me: and the sea saith, It uentium 28:14 abyssus dicit non est in me et mare loquitur is not with me. 28:15 It cannot be gotten for gold, neither shall silver be non est mecum 28:15 non dabitur aurum obrizum pro ea weighed for the price thereof. 28:16 It cannot be valued with the gold of Ophir, with the nec adpendetur argentum in commutatione eius 28:16 non precious onyx, or the sapphire.

conferetur tinctis Indiae coloribus nec lapidi sardonico pre-

28:17 The gold and the crystal cannot equal it: and the tiosissimo uel sapphyro 28:17 non adaequabitur ei aurum

exchange of it shall not be for jewels of fine gold. 28:18 No mention shall be made of coral, or of pearls: for uel uitrum nec commutabuntur pro ea uasa auri 28:18 exthe price of wisdom is above rubies.

celsa et eminentia non memorabuntur conparatione eius tra-

28:19 The topaz of Ethiopia shall not equal it, neither shall hitur autem sapientia de occultis 28:19 non adaequabitur ei it be valued with pure gold.

topazium de Aethiopia nec tincturae mundissimae conpone-

28:20 Whence then cometh wisdom? and where is the tur 28:20 unde ergo sapientia ueniet et quis est locus intel-

place of understanding?

28:21 Seeing it is hid from the eyes of all living, and kept legentiae 28:21 abscondita est ab oculis omnium uiuentium

close from the fowls of the air.

28:22 Destruction and death say, We have heard the fame uolucres quoque caeli latet 28:22 perditio et mors dixerunt

thereof with our ears.

28:23 God understandeth the way thereof, and he knoweth auribus nostris audiuimus famam eius 28:23 Deus intellegit

the place thereof.

uiam eius et ipse nouit locum illius 28:24 ipse enim fines 28:24 For he looketh to the ends of the earth, and seeth

under the whole heaven;

mundi intuetur et omnia quae sub caelo sunt respicit 28:25 28:25 To make the weight for the winds; and he weigheth

the waters by measure.

qui fecit uentis pondus et aquas adpendit mensura 28:26 28:26 When he made a decree for the rain, and a way for

the lightning of the thunder:

quando ponebat pluuiis legem et uiam procellis sonantibus 28:27 tunc uidit illam et enarrauit et praeparauit et inuestiga- 28:27 Then did he see it, and declare it; he prepared it,

yea, and searched it out.

uit 28:28 et dixit homini ecce timor Domini ipsa est sapien- 28:28 And unto man he said, Behold, the fear of the

LORD, that is wisdom; and to depart from evil is understanding.

tia et recedere a malo intellegentia

29:1 addidit quoque Iob adsumens parabolam suam et dixit 29:1 Moreover Job continued his parable, and said, 29:2 quis mihi tribuat ut sim iuxta menses pristinos secundum 29:2 Oh that I were as in months past, as in the days when

God preserved me;

dies quibus Deus custodiebat me 29:3 quando splendebat 29:3 When his candle shined upon my head, and when by

his light I walked through darkness;

lucerna eius super caput meum et ad lumen eius ambulabam in tenebris 29:4 sicut fui in diebus adulescentiae meae 29:4 As I was in the days of my youth, when the secret of

God was upon my tabernacle;

quando secreto Deus erat in tabernaculo meo 29:5 quando 29:5 When the Almighty was yet with me, when my chil-

dren were about me;

erat Omnipotens mecum et in circuitu meo pueri mei 29:6 29:6 When I washed my steps with butter, and the rock

poured me out rivers of oil;

quando lauabam pedes meos butyro et petra fundebat mihi riuos olei 29:7 quando procedebam ad portam ciuitatis et 29:7 When I went out to the gate through the city, when I

prepared my seat in the street!

in platea parabant cathedram mihi 29:8 uidebant me iu- 29:8 The young men saw me, and hid themselves: and the

aged arose, and stood up.

uenes et abscondebantur et senes adsurgentes stabant 29:9 29:9 The princes refrained talking, and laid their hand on

their mouth.

principes cessabant loqui et digitum superponebant ori suo 29:10 uocem suam cohibebant duces et lingua eorum gutturi 29:10 The nobles held their peace, and their tongue

cleaved to the roof of their mouth.

suo adherebat 29:11 auris audiens beatificabat me et oculus 29:11 When the ear heard me, then it blessed me; and

when the eye saw me, it gave witness to me:

uidens testimonium reddebat mihi 29:12 quod liberassem 29:12 Because I delivered the poor that cried, and the fa-

therless, and him that had none to help him.

pauperem uociferantem et pupillum cui non esset adiutor 29:13 benedictio perituri super me ueniebat et cor uiduae con- 29:13 The blessing of him that was ready to perish came

upon me: and I caused the widow’s heart to sing for joy.

solatus sum 29:14 iustitia indutus sum et uestiuit me sic- 29:14 I put on righteousness, and it clothed me: my judg-

ment was as a robe and a diadem.

ut uestimento et diademate iudicio meo 29:15 oculus fui 29:15 I was eyes to the blind, and feet was I to the lame. caeco et pes claudo 29:16 pater eram pauperum et causam 29:16 I was a father to the poor: and the cause which I

knew not I searched out.

quam nesciebam diligentissime inuestigabam 29:17 conte- 29:17 And I brake the jaws of the wicked, and plucked the

spoil out of his teeth.

rebam molas iniqui et de dentibus illius auferebam praedam 29:18 dicebamque in nidulo meo moriar et sicut palma mul- 29:18 Then I said, I shall die in my nest, and I shall multi-

ply my days as the sand.

tiplicabo dies 29:19 radix mea aperta est secus aquas et ros 29:19 My root was spread out by the waters, and the dew

lay all night upon my branch.

morabitur in messione mea 29:20 gloria mea semper inno- 29:20 My glory was fresh in me, and my bow was renewed

in my hand.

uabitur et arcus meus in manu mea instaurabitur 29:21 qui 29:21 Unto me men gave ear, and waited, and kept silence

at my counsel.

me audiebant expectabant sententiam et intenti tacebant ad

29:22 After my words they spake not again; and my consilium meum 29:22 uerbis meis addere nihil audebant et

speech dropped upon them.

29:23 And they waited for me as for the rain; and they super illos stillabat eloquium meum 29:23 expectabant me

opened their mouth wide as for the latter rain.

sicut pluuiam et os suum aperiebant quasi ad imbrem sero-

29:24 If I laughed on them, they believed it not; and the tinum 29:24 si quando ridebam ad eos non credebant et lux

light of my countenance they cast not down.

29:25 I chose out their way, and sat chief, and dwelt as a uultus mei non cadebat in terram 29:25 si uoluissem ire ad

king in the army, as one that comforteth the mourners.

eos sedebam primus cumque sederem quasi rex circumstante exercitu eram tamen maerentium consolator

30:1 But now they that are younger than I have me in de- 30:1 nunc autem derident me iuniores tempore quorum rision, whose fathers I would have disdained to have set with the dogs of my flock.

non dignabar patres ponere cum canibus gregis mei 30:2

30:2 Yea, whereto might the strength of their hands profit me, in whom old age was perished?

quorum uirtus manuum erat mihi pro nihilo et uita ipsa pu-

30:3 For want and famine they were solitary; fleeing into tabantur indigni 30:3 egestate et fame steriles qui rodebant

the wilderness in former time desolate and waste.

30:4 Who cut up mallows by the bushes, and juniper roots in solitudine squalentes calamitate et miseria 30:4 et man-

for their meat.

debant herbas et arborum cortices et radix iuniperorum erat

30:5 They were driven forth from among men, (they cried cibus eorum 30:5 qui de conuallibus ista rapientes cum sin-

after them as after a thief;)

30:6 To dwell in the cliffs of the valleys, in caves of the gula repperissent ad ea cum clamore currebant 30:6 in de-

earth, and in the rocks.

sertis habitabant torrentium et in cauernis terrae uel super

30:7 Among the bushes they brayed; under the nettles they glaream 30:7 qui inter huiuscemodi laetabantur et esse sub

were gathered together.

30:8 They were children of fools, yea, children of base sentibus delicias conputabant 30:8 filii stultorum et ignobi-

men: they were viler than the earth.

30:9 And now am I their song, yea, I am their byword. lium et in terra penitus non parentes 30:9 nunc in eorum 30:10 They abhor me, they flee far from me, and spare not canticum uersus sum et factus sum eis prouerbium 30:10

to spit in my face.

abominantur me et longe fugiunt a me et faciem meam con-

30:11 Because he hath loosed my cord, and afflicted me, spuere non uerentur 30:11 faretram enim suam aperuit et

they have also let loose the bridle before me.

30:12 Upon my right hand rise the youth; they push away adflixit me et frenum posuit in os meum 30:12 ad dexteram

my feet, and they raise up against me the ways of their destruction.

orientis calamitatis meae ilico surrexerunt pedes meos sub-

30:13 They mar my path, they set forward my calamity, uerterunt et oppresserunt quasi fluctibus semitis suis 30:13

they have no helper.

dissipauerunt itinera mea insidiati sunt mihi et praeualue-

30:14 They came upon me as a wide breaking in of waters: runt et non fuit qui ferret auxilium 30:14 quasi rupto muro

in the desolation they rolled themselves upon me.

et aperta ianua inruerunt super me et ad meas miserias deuo-

30:15 Terrors are turned upon me: they pursue my soul as luti sunt 30:15 redactus sum in nihili abstulisti quasi uentus

the wind: and my welfare passeth away as a cloud.

30:16 And now my soul is poured out upon me; the days desiderium meum et uelut nubes pertransiit salus mea 30:16

of affliction have taken hold upon me.

nunc autem in memet ipso marcescit anima mea et possident

30:17 My bones are pierced in me in the night season: and me dies adflictionis 30:17 nocte os meum perforatur dolo-

my sinews take no rest.

ribus et qui me comedunt non dormiunt 30:18 in multitu- 30:18 By the great force of my disease is my garment

changed: it bindeth me about as the collar of my coat.

dine eorum consumitur uestimentum meum et quasi capitio tunicae sic cinxerunt me 30:19 conparatus sum luto et adsi- 30:19 He hath cast me into the mire, and I am become like

dust and ashes.

milatus fauillae et cineri 30:20 clamo ad te et non exaudis 30:20 I cry unto thee, and thou dost not hear me: I stand

up, and thou regardest me not.

me sto et non respicis me 30:21 mutatus es mihi in crude- 30:21 Thou art become cruel to me: with thy strong hand

thou opposest thyself against me.

lem et in duritia manus tuae aduersaris mihi 30:22 eleuasti 30:22 Thou liftest me up to the wind; thou causest me to

ride upon it, and dissolvest my substance.

me et quasi super uentum ponens elisisti me ualide 30:23 30:23 For I know that thou wilt bring me to death, and to

the house appointed for all living.

scio quia morti tradas me ubi constituta domus est omni

uiuenti 30:24 uerumtamen non ad consumptionem eorum 30:24 Howbeit he will not stretch out his hand to the grave,

though they cry in his destruction.

emittis manum tuam et si corruerint ipse saluabis 30:25 fle- 30:25 Did not I weep for him that was in trouble? was not

my soul grieved for the poor?

bam quondam super eum qui adflictus erat et conpatiebatur anima mea pauperi 30:26 expectabam bona et uenerunt 30:26 When I looked for good, then evil came unto me:

and when I waited for light, there came darkness.

mihi mala praestolabar lucem et eruperunt tenebrae 30:27 30:27 My bowels boiled, and rested not: the days of afflic-

tion prevented me.

interiora mea efferbuerunt absque ulla requie praeuenerunt me dies adflictionis 30:28 maerens incedebam sine furore 30:28 I went mourning without the sun: I stood up, and I

cried in the congregation.

consurgens in turba clamaui 30:29 frater fui draconum et 30:29 I am a brother to dragons, and a companion to owls. socius strutionum 30:30 cutis mea denigrata est super me 30:30 My skin is black upon me, and my bones are burned

with heat.

et ossa mea aruerunt prae caumate 30:31 uersa est in luc- 30:31 My harp also is turned to mourning, and my organ

into the voice of them that weep.

tum cithara mea et organum meum in uocem flentium

31:1 pepigi foedus cum oculis meis ut ne cogitarem qui- 31:1 I made a covenant with mine eyes; why then should I

think upon a maid?

dem de uirgine 31:2 quam enim partem haberet Deus in me 31:2 For what portion of God is there from above? and

what inheritance of the Almighty from on high?

desuper et hereditatem Omnipotens de excelsis 31:3 num- 31:3 Is not destruction to the wicked? and a strange pun-

ishment to the workers of iniquity?

quid non perditio est iniquo et alienatio operantibus iniustitiam 31:4 nonne ipse considerat uias meas et cunctos gres- 31:4 Doth not he see my ways, and count all my steps? sus meos dinumerat 31:5 si ambulaui in uanitate et festi- 31:5 If I have walked with vanity, or if my foot hath hasted

to deceit;

nauit in dolo pes meus 31:6 adpendat me in statera iusta 31:6 Let me be weighed in an even balance that God may

know mine integrity.

et sciat Deus simplicitatem meam 31:7 si declinauit gres- 31:7 If my step hath turned out of the way, and mine heart

walked after mine eyes, and if any blot hath cleaved to mine hands;

sus meus de uia et si secutum est oculos meos cor meum et in manibus meis adhesit macula 31:8 seram et alius co- 31:8 Then let me sow, and let another eat; yea, let my

offspring be rooted out.

medat et progenies mea eradicetur 31:9 si deceptum est cor 31:9 If mine heart have been deceived by a woman, or if I

have laid wait at my neighbour’s door;

meum super mulierem et si ad ostium amici mei insidiatus sum 31:10 scortum sit alteri uxor mea et super illam incur- 31:10 Then let my wife grind unto another, and let others

bow down upon her.

uentur alii 31:11 hoc enim nefas est et iniquitas maxima 31:11 For this is an heinous crime; yea, it is an iniquity to be punished by the judges.

31:12 For it is a fire that consumeth to destruction, and 31:12 ignis est usque ad perditionem deuorans et omnia era-

would root out all mine increase.

31:13 If I did despise the cause of my manservant or of my dicans genimina 31:13 si contempsi subire iudicium cum

maidservant, when they contended with me;

seruo meo et ancillae meae cum disceptarent aduersum me

31:14 What then shall I do when God riseth up? and when 31:14 quid enim faciam cum surrexerit ad iudicandum Deus

he visiteth, what shall I answer him?

31:15 Did not he that made me in the womb make him? et cum quaesierit quid respondebo illi 31:15 numquid non

and did not one fashion us in the womb?

in utero fecit me qui et illum operatus est et formauit in uulua

31:16 If I have withheld the poor from their desire, or have unus 31:16 si negaui quod uolebant pauperibus et oculos ui-

caused the eyes of the widow to fail;

31:17 Or have eaten my morsel myself alone, and the fa- duae expectare feci 31:17 si comedi buccellam meam solus

therless hath not eaten thereof;

31:18 (For from my youth he was brought up with me, et non comedit pupillus ex ea 31:18 quia ab infantia mea

as with a father, and I have guided her from my mother’s womb;)

creuit mecum miseratio et de utero matris meae egressa est

31:19 If I have seen any perish for want of clothing, or any mecum 31:19 si despexi pereuntem eo quod non habuerit poor without covering; 31:20 If his loins have not blessed me, and if he were not indumentum et absque operimento pauperem 31:20 si non warmed with the fleece of my sheep;

benedixerunt mihi latera eius et de uelleribus ouium mea-

31:21 If I have lifted up my hand against the fatherless, rum calefactus est 31:21 si leuaui super pupillum manum when I saw my help in the gate: 31:22 Then let mine arm fall from my shoulder blade, and meam etiam cum uiderem me in porta superiorem 31:22 mine arm be broken from the bone.

umerus meus a iunctura sua cadat et brachium meum cum

31:23 For destruction from God was a terror to me, and by suis ossibus confringatur 31:23 semper enim quasi tumen-

reason of his highness I could not endure.

tes super me fluctus timui Deum et pondus eius ferre non

31:24 If I have made gold my hope, or have said to the fine potui 31:24 si putaui aurum robur meum et obrizae dixi fi-

gold, Thou art my confidence;

31:25 If I rejoice because my wealth was great, and be- ducia mea 31:25 si laetatus sum super multis diuitiis meis

cause mine hand had gotten much;

31:26 If I beheld the sun when it shined, or the moon walk- et quia plurima repperit manus mea 31:26 si uidi solem cum

ing in brightness;

31:27 And my heart hath been secretly enticed, or my fulgeret et lunam incedentem clare 31:27 et lactatum est in

mouth hath kissed my hand:

abscondito cor meum et osculatus sum manum meam ore

31:28 This also were an iniquity to be punished by the meo 31:28 quae est iniquitas maxima et negatio contra De-

judge: for I should have denied the God that is above. 31:29 If I rejoice at the destruction of him that hated me, um altissimum 31:29 si gauisus sum ad ruinam eius qui me or lifted up myself when evil found him:

31:30 Neither have I suffered my mouth to sin by wishing oderat et exultaui quod inuenisset eum malum 31:30 non

a curse to his soul.

enim dedi ad peccandum guttur meum ut expeterem maledi-

31:31 If the men of my tabernacle said not, Oh that we had cens animam eius 31:31 si non dixerunt uiri tabernaculi mei

of his flesh! we cannot be satisfied. 31:32 The stranger did not lodge in the street: but I opened quis det de carnibus eius ut saturemur 31:32 foris non manmy doors to the traveller.

31:33 If I covered my transgressions as Adam, by hiding sit peregrinus ostium meum uiatori patuit 31:33 si abscondi

mine iniquity in my bosom:

quasi homo peccatum meum et celaui in sinu meo iniquita-

31:34 Did I fear a great multitude, or did the contempt of tem meam 31:34 si expaui ad multitudinem nimiam et defamilies terrify me, that I kept silence, and went not out of the door?

spectio propinquorum terruit me et non magis tacui nec egressus sum ostium 31:35 quis mihi tribuat auditorem ut de- 31:35 Oh that one would hear me! behold, my desire is,

that the Almighty would answer me, and that mine adversary had written a book.

siderium meum Omnipotens audiat et librum scribat ipse qui iudicat 31:36 ut in umero meo portem illum et circumdem 31:36 Surely I would take it upon my shoulder, and bind

it as a crown to me.

illum quasi coronam mihi 31:37 per singulos gradus meos 31:37 I would declare unto him the number of my steps;

as a prince would I go near unto him.

pronuntiabo illum et quasi principi offeram eum 31:38 si 31:38 If my land cry against me, or that the furrows like-

wise thereof complain;

aduersum me terra mea clamat et cum ipsa sulci eius deflent 31:39 si fructus eius comedi absque pecunia et animam agri- 31:39 If I have eaten the fruits thereof without money, or

have caused the owners thereof to lose their life:

colarum eius adflixi 31:40 pro frumento oriatur mihi tribu- 31:40 Let thistles grow instead of wheat, and cockle in-

stead of barley. The words of Job are ended.

lus et pro hordeo spina finita sunt uerba Iob

32:1 omiserunt autem tres uiri isti respondere Iob eo quod 32:1 So these three men ceased to answer Job, because he

was righteous in his own eyes.

iustus sibi uideretur 32:2 et iratus indignatusque Heliu fi- 32:2 Then was kindled the wrath of Elihu the son of

Barachel the Buzite, of the kindred of Ram: against Job was his wrath kindled, because he justified himself rather

lius Barachel Buzites de cognatione Ram iratus est autem than God. aduersus Iob eo quod iustum se esse diceret coram Deo 32:3 porro aduersum amicos eius indignatus est eo quod 32:3 Also against his three friends was his wrath kin-

dled, because they had found no answer, and yet had condemned Job.

non inuenissent responsionem rationabilem sed tantummodo condemnassent Iob 32:4 igitur Heliu expectauit Iob loquen- 32:4 Now Elihu had waited till Job had spoken, because

they were elder than he.

tem eo quod seniores se essent qui loquebantur 32:5 cum 32:5 When Elihu saw that there was no answer in the

mouth of these three men, then his wrath was kindled.

autem uidisset quod tres respondere non potuissent iratus est uehementer 32:6 respondensque Heliu filius Barachel 32:6 And Elihu the son of Barachel the Buzite answered

and said, I am young, and ye are very old; wherefore I was afraid, and durst not shew you mine opinion.

Buzites dixit iunior sum tempore uos autem antiquiores idcirco dimisso capite ueritus sum indicare uobis meam sententiam 32:7 sperabam enim quod aetas prolixior loquere- 32:7 I said, Days should speak, and multitude of years

should teach wisdom.

tur et annorum multitudo doceret sapientiam 32:8 sed ut ui- 32:8 But there is a spirit in man: and the inspiration of the

Almighty giveth them understanding.

deo spiritus est in hominibus et inspiratio Omnipotentis dat intellegentiam 32:9 non sunt longeui sapientes nec senes 32:9 Great men are not always wise: neither do the aged

understand judgment.

intellegunt iudicium 32:10 ideo dicam audite me ostendam 32:10 Therefore I said, Hearken to me; I also will shew

mine opinion.

uobis etiam ego meam scientiam 32:11 expectaui enim ser- 32:11 Behold, I waited for your words; I gave ear to your

reasons, whilst ye searched out what to say.

mones uestros audiui prudentiam uestram donec disceptaremini sermonibus 32:12 et donec putabam uos aliquid dicere 32:12 Yea, I attended unto you, and, behold, there was

none of you that convinced Job, or that answered his words:

considerabam sed ut uideo non est qui arguere possit Iob et respondere ex uobis sermonibus eius 32:13 ne forte dicatis 32:13 Lest ye should say, We have found out wisdom: God

thrusteth him down, not man.

inuenimus sapientiam Deus proiecit eum non homo 32:14 32:14 Now he hath not directed his words against me: nei-

ther will I answer him with your speeches.

nihil locutus est mihi et ego non secundum uestros sermones 32:15 They were amazed, they answered no more: they respondebo illi 32:15 extimuerunt non responderunt ultra

left off speaking.

32:16 When I had waited, (for they spake not, but stood abstuleruntque a se eloquia 32:16 quoniam igitur expectaui

still, and answered no more;)

32:17 I said, I will answer also my part, I also will shew et non sunt locuti steterunt nec responderunt ultra 32:17 re-

mine opinion.

spondebo et ego partem meam et ostendam scientiam meam

32:18 For I am full of matter, the spirit within me con- 32:18 plenus sum enim sermonibus et coartat me spiritus uteri

straineth me.

32:19 Behold, my belly is as wine which hath no vent; it mei 32:19 en uenter meus quasi mustum absque spiraculo

is ready to burst like new bottles.

32:20 I will speak, that I may be refreshed: I will open my quod lagunculas nouas disrumpit 32:20 loquar et respirabo

lips and answer.

32:21 Let me not, I pray you, accept any man’s person, paululum aperiam labia mea et respondebo 32:21 non ac-

neither let me give flattering titles unto man. 32:22 For I know not to give flattering titles; in so doing cipiam personam uiri et Deum homini non aequabo 32:22 my maker would soon take me away.

nescio enim quamdiu subsistam et si post modicum tollat me factor meus

33:1 Wherefore, Job, I pray thee, hear my speeches, and 33:1 audi igitur Iob eloquia mea et omnes sermones meos hearken to all my words.

33:2 Behold, now I have opened my mouth, my tongue ausculta 33:2 ecce aperui os meum loquatur lingua mea

hath spoken in my mouth.

33:3 My words shall be of the uprightness of my heart: in faucibus meis 33:3 simplici corde meo sermones mei et

and my lips shall utter knowledge clearly.

33:4 The spirit of God hath made me, and the breath of the sententiam labia mea puram loquentur 33:4 spiritus Dei fe-

Almighty hath given me life. 33:5 If thou canst answer me, set thy words in order before cit me et spiraculum Omnipotentis uiuificauit me 33:5 si me, stand up.

33:6 Behold, I am according to thy wish in God’s stead: I potes responde mihi et aduersus faciem meam consiste 33:6

also am formed out of the clay.

ecce et me sicut et te fecit Deus et de eodem luto ego quo-

33:7 Behold, my terror shall not make thee afraid, neither que formatus sum 33:7 uerumtamen miraculum meum non shall my hand be heavy upon thee.

33:8 Surely thou hast spoken in mine hearing, and I have te terreat et eloquentia mea non sit tibi grauis 33:8 dix-

heard the voice of thy words, saying, 33:9 I am clean without transgression, I am innocent; nei- isti ergo in auribus meis et uocem uerborum audiui 33:9 ther is there iniquity in me.

mundus sum ego absque delicto inmaculatus et non est in-

33:10 Behold, he findeth occasions against me, he coun- iquitas in me 33:10 quia querellas in me repperit ideo ar-

teth me for his enemy,

33:11 He putteth my feet in the stocks, he marketh all my bitratus est me inimicum sibi 33:11 posuit in neruo pedes

paths. 33:12 Behold, in this thou art not just: I will answer thee, meos custodiuit omnes semitas meas 33:12 hoc est ergo in that God is greater than man.

quo non es iustificatus respondebo tibi quia maior sit Deus

33:13 Why dost thou strive against him? for he giveth not homine 33:13 aduersum eum contendis quod non ad omaccount of any of his matters.

33:14 For God speaketh once, yea twice, yet man per- nia uerba responderit tibi 33:14 semel loquitur Deus et se-

ceiveth it not. 33:15 In a dream, in a vision of the night, when deep sleep cundo id ipsum non repetit 33:15 per somnium in uisione falleth upon men, in slumberings upon the bed;

nocturna quando inruit sopor super homines et dormiunt in

33:16 Then he openeth the ears of men, and sealeth their lectulo 33:16 tunc aperit aures uirorum et erudiens eos in-

instruction,

33:17 That he may withdraw man from his purpose, and struit disciplinam 33:17 ut auertat hominem ab his quae fa-

hide pride from man.

33:18 He keepeth back his soul from the pit, and his life cit et liberet eum de superbia 33:18 eruens animam eius a

from perishing by the sword.

corruptione et uitam illius ut non transeat in gladium 33:19 33:19 He is chastened also with pain upon his bed, and the

multitude of his bones with strong pain:

increpat quoque per dolorem in lectulo et omnia ossa eius marcescere facit 33:20 abominabilis ei fit in uita sua panis 33:20 So that his life abhorreth bread, and his soul dainty

meat.

et animae illius cibus ante desiderabilis 33:21 tabescet caro 33:21 His flesh is consumed away, that it cannot be seen;

and his bones that were not seen stick out.

eius et ossa quae tecta fuerant nudabuntur 33:22 adpropin- 33:22 Yea, his soul draweth near unto the grave, and his

life to the destroyers.

quabit corruptioni anima eius et uita illius mortiferis 33:23 33:23 If there be a messenger with him, an interpreter, one

among a thousand, to shew unto man his uprightness:

si fuerit pro eo angelus loquens unum de milibus ut adnuntiet hominis aequitatem 33:24 miserebitur eius et dicet li- 33:24 Then he is gracious unto him, and saith, Deliver him

from going down to the pit: I have found a ransom.

bera eum et non descendat in corruptionem inueni in quo ei propitier 33:25 consumpta est caro eius a suppliciis reuer- 33:25 His flesh shall be fresher than a child’s: he shall

return to the days of his youth:

tatur ad dies adulescentiae suae 33:26 deprecabitur Deum 33:26 He shall pray unto God, and he will be favourable

unto him: and he shall see his face with joy: for he will render unto man his righteousness.

et placabilis ei erit et uidebit faciem eius in iubilo et reddet homini iustitiam suam 33:27 respiciet homines et dicet 33:27 He looketh upon men, and if any say, I have sinned,

and perverted that which was right, and it profited me not;

peccaui et uere deliqui et ut eram dignus non recepi 33:28 33:28 He will deliver his soul from going into the pit, and

his life shall see the light.

liberauit animam suam ne pergeret in interitum sed uiuens lucem uideret 33:29 ecce haec omnia operatur Deus tribus 33:29 Lo, all these things worketh God oftentimes with

man,

uicibus per singulos 33:30 ut reuocet animas eorum a cor- 33:30 To bring back his soul from the pit, to be enlightened

with the light of the living.

ruptione et inluminet luce uiuentium 33:31 adtende Iob et 33:31 Mark well, O Job, hearken unto me: hold thy peace,

and I will speak.

audi me et tace dum ego loquar 33:32 si autem habes quod 33:32 If thou hast anything to say, answer me: speak, for

I desire to justify thee.

loquaris responde mihi loquere uolo enim te apparere iustum 33:33 quod si non habes audi me tace et docebo te sapientiam 33:33 If not, hearken unto me: hold thy peace, and I shall

teach thee wisdom.

34:1 pronuntians itaque Heliu etiam haec locutus est 34:2 34:1 Furthermore Elihu answered and said,

34:2 Hear my words, O ye wise men; and give ear unto me, ye that have knowledge.

audite sapientes uerba mea et eruditi auscultate me 34:3 au-

34:3 For the ear trieth words, as the mouth tasteth meat.

ris enim uerba probat et guttur escas gustu diiudicat 34:4 34:4 Let us choose to us judgment: let us know among

ourselves what is good.

iudicium eligamus nobis et inter nos uideamus quid sit melius 34:5 quia dixit Iob iustus sum et Deus subuertit iudi- 34:5 For Job hath said, I am righteous: and God hath taken

away my judgment.

cium meum 34:6 in iudicando enim me mendacium est uio- 34:6 Should I lie against my right? my wound is incurable

without transgression.

lenta sagitta mea absque ullo peccato 34:7 quis est uir ut est 34:7 What man is like Job, who drinketh up scorning like

water?

Iob qui bibit subsannationem quasi aquam 34:8 qui gradi- 34:8 Which goeth in company with the workers of iniq-

uity, and walketh with wicked men.

tur cum operantibus iniquitatem et ambulat cum uiris impiis 34:9 dixit enim non placebit uir Deo etiam si cucurrerit cum 34:9 For he hath said, It profiteth a man nothing that he

should delight himself with God.

eo 34:10 ideo uiri cordati audite me absit a Deo impietas 34:10 Therefore hearken unto me ye men of understand-

ing: far be it from God, that he should do wickedness; and from the Almighty, that he should commit iniquity.

et ab Omnipotente iniquitas 34:11 opus enim hominis re-

34:11 For the work of a man shall he render unto him, and cause every man to find according to his ways.

34:12 Yea, surely God will not do wickedly, neither will ddet ei et iuxta uias singulorum restituet 34:12 uere enim

the Almighty pervert judgment.

Deus non condemnabit frustra nec Omnipotens subuertet iu-

34:13 Who hath given him a charge over the earth? or who dicium 34:13 quem constituit alium super terram aut quem

hath disposed the whole world?

34:14 If he set his heart upon man, if he gather unto him- posuit super orbem quem fabricatus est 34:14 si direxerit ad

self his spirit and his breath;

34:15 All flesh shall perish together, and man shall turn eum cor suum spiritum illius et flatum ad se trahet 34:15 de-

again unto dust.

34:16 If now thou hast understanding, hear this: hearken ficiet omnis caro simul et homo in cinerem reuertetur 34:16

to the voice of my words.

si habes ergo intellectum audi quod dicitur et ausculta uo-

34:17 Shall even he that hateth right govern? and wilt thou cem eloquii mei 34:17 numquid qui non amat iudicium sacondemn him that is most just?

nare potest et quomodo tu eum qui iustus est in tantum con-

34:18 Is it fit to say to a king, Thou art wicked? and to demnas 34:18 qui dicit regi apostata qui uocat duces im-

princes, Ye are ungodly?

34:19 How much less to him that accepteth not the persons pios 34:19 qui non accipit personas principum nec cognouit

of princes, nor regardeth the rich more than the poor? for they all are the work of his hands.

tyrannum cum disceptaret contra pauperem opus enim ma-

34:20 In a moment shall they die, and the people shall be nuum eius sunt uniuersi 34:20 subito morientur et in media

troubled at midnight, and pass away: and the mighty shall be taken away without hand.

nocte turbabuntur populi et pertransibunt et auferent uiolen-

34:21 For his eyes are upon the ways of man, and he seeth tum absque manu 34:21 oculi enim eius super uias homiall his goings.

34:22 There is no darkness, nor shadow of death, where num et omnes gressus eorum considerat 34:22 non sunt te-

the workers of iniquity may hide themselves.

nebrae et non est umbra mortis ut abscondantur ibi qui ope-

34:23 For he will not lay upon man more than right; that rantur iniquitatem 34:23 neque enim ultra in hominis pote-

he should enter into judgment with God.

34:24 He shall break in pieces mighty men without num- state est ut ueniat ad Deum in iudicium 34:24 conteret mul-

ber, and set others in their stead.

34:25 Therefore he knoweth their works, and he overtur- tos innumerabiles et stare faciet alios pro eis 34:25 nouit

neth them in the night, so that they are destroyed.

enim opera eorum et idcirco inducet noctem et conterentur

34:26 He striketh them as wicked men in the open sight of 34:26 quasi impios percussit eos in loco uidentium 34:27 qui

others; 34:27 Because they turned back from him, and would not consider any of his ways:

quasi de industria recesserunt ab eo et omnes uias eius in-

34:28 So that they cause the cry of the poor to come unto tellegere noluerunt 34:28 ut peruenire facerent ad eum cla-

him, and he heareth the cry of the afflicted. 34:29 When he giveth quietness, who then can make trou- morem egeni et audiret uocem pauperum 34:29 ipso enim ble? and when he hideth his face, who then can behold him? whether it be done against a nation, or against a man

only: concedente pacem quis est qui condemnet ex quo absconderit uultum quis est qui contempletur eum et super gentem et

34:30 That the hypocrite reign not, lest the people be en- super omnes homines 34:30 qui regnare facit hominem hysnared.

34:31 Surely it is meet to be said unto God, I have borne pocritam propter peccata populi 34:31 quia ergo ego locu-

chastisement, I will not offend any more:

34:32 That which I see not teach thou me: if I have done tus sum ad Deum te quoque non prohibeo 34:32 si erraui tu

iniquity, I will do no more.

34:33 Should it be according to thy mind? he will recom- doce me si iniquitatem locutus sum ultra non addam 34:33

pense it, whether thou refuse, or whether thou choose; and not I: therefore speak what thou knowest.

numquid a te Deus expetit eam quia displicuit tibi tu enim coepisti loqui et non ego quod si quid nosti melius loquere 34:34 uiri intellegentes loquantur mihi et uir sapiens audiat 34:34 Let men of understanding tell me, and let a wise

man hearken unto me.

me 34:35 Iob autem stulte locutus est et uerba illius non 34:35 Job hath spoken without knowledge, and his words

were without wisdom.

sonant disciplinam 34:36 pater mi probetur Iob usque ad 34:36 My desire is that Job may be tried unto the end be-

cause of his answers for wicked men.

finem ne desinas in hominibus iniquitatis 34:37 quia addit 34:37 For he addeth rebellion unto his sin, he clappeth his

hands among us, and multiplieth his words against God.

super peccata sua blasphemiam inter nos interim constringatur et tunc ad iudicium prouocet sermonibus suis Deum

35:1 igitur Heliu haec rursum locutus est 35:2 numquid 35:1 Elihu spake moreover, and said,

35:2 Thinkest thou this to be right, that thou saidst, My righteousness is more than God’s?

aequa tibi uidetur tua cogitatio ut diceres iustior Deo sum 35:3 dixisti enim non tibi placet quod rectum est uel quid 35:3 For thou saidst, What advantage will it be unto thee?

and, What profit shall I have, if I be cleansed from my sin?

tibi prodeeit si ego peccauero 35:4 itaque ego respondebo 35:4 I will answer thee, and thy companions with thee. sermonibus tuis et amicis tuis tecum 35:5 suspice caelum 35:5 Look unto the heavens, and see; and behold the

clouds which are higher than thou.

et intuere et contemplare aethera quod altior te sit 35:6 si 35:6 If thou sinnest, what doest thou against him? or if thy

transgressions be multiplied, what doest thou unto him?

peccaueris quid ei nocebis et si multiplicatae fuerint iniquitates tuae quid facies contra eum 35:7 porro si iuste egeris 35:7 If thou be righteous, what givest thou him? or what

receiveth he of thine hand?

quid donabis ei aut quid de manu tua accipiet 35:8 homini 35:8 Thy wickedness may hurt a man as thou art; and thy

righteousness may profit the son of man.

qui similis tui est nocebit impietas tua et filium hominis adiuuabit iustitia tua 35:9 propter multitudinem calumniato- 35:9 By reason of the multitude of oppressions they make

the oppressed to cry: they cry out by reason of the arm of the mighty.

rum clamabunt et heiulabunt propter uim brachii tyranno-

rum 35:10 et non dixit ubi est Deus qui fecit me qui de- 35:10 But none saith, Where is God my maker, who giveth

songs in the night;

dit carmina in nocte 35:11 qui docet nos super iumenta ter- 35:11 Who teacheth us more than the beasts of the earth,

and maketh us wiser than the fowls of heaven?

rae et super uolucres caeli erudit nos 35:12 ibi clamabunt 35:12 There they cry, but none giveth answer, because of

the pride of evil men.

et non exaudiet propter superbiam malorum 35:13 non ergo 35:13 Surely God will not hear vanity, neither will the

Almighty regard it.

frustra audiet Deus et Omnipotens singulorum causas intuebitur 35:14 etiam cum dixeris non considerat iudicare co- 35:14 Although thou sayest thou shalt not see him, yet

judgment is before him; therefore trust thou in him.

ram eo et expecta eum 35:15 nunc enim non infert furorem 35:15 But now, because it is not so, he hath visited in his

anger; yet he knoweth it not in great extremity:

suum nec ulciscitur scelus ualde 35:16 ergo Iob frustra ape- 35:16 Therefore doth Job open his mouth in vain; he mul-

tiplieth words without knowledge.

rit os suum et absque scientia uerba multiplicat

36:1 addens quoque Heliu haec locutus est 36:2 sustine 36:1 Elihu also proceeded, and said,

36:2 Suffer me a little, and I will shew thee that I have yet to speak on God’s behalf.

me paululum et indicabo tibi adhuc enim habeo quod pro

Deo loquar 36:3 repetam scientiam meam a principio et 36:3 I will fetch my knowledge from afar, and will ascribe

righteousness to my Maker.

operatorem meum probabo iustum 36:4 uere enim absque 36:4 For truly my words shall not be false: he that is per-

fect in knowledge is with thee.

mendacio sermones mei et perfecta scientia probabitur tibi 36:5 Deus potentes non abicit cum et ipse sit potens 36:6 36:5 Behold, God is mighty, and despiseth not any: he is

mighty in strength and wisdom.

36:6 He preserveth not the life of the wicked: but giveth right to the poor.

36:7 He withdraweth not his eyes from the righteous: but sed non saluat impios et iudicium pauperibus tribuit 36:7

with kings are they on the throne; yea, he doth establish them for ever, and they are exalted.

non aufert a iusto oculos suos et reges in solio conlocat in

36:8 And if they be bound in fetters, and be holden in perpetuum et illi eriguntur 36:8 et si fuerint in catenis et

cords of affliction;

36:9 Then he sheweth them their work, and their trans- uinciantur funibus paupertatis 36:9 indicabit eis opera eo-

gressions that they have exceeded.

36:10 He openeth also their ear to discipline, and com- rum et scelera eorum quia uiolenti fuerint 36:10 reuelabit

mandeth that they return from iniquity.

quoque aurem eorum ut corripiat et loquetur ut reuertantur

36:11 If they obey and serve him, they shall spend their ab iniquitate 36:11 si audierint et obseruauerint conplebunt

days in prosperity, and their years in pleasures. 36:12 But if they obey not, they shall perish by the sword, dies suos in bono et annos suos in gloria 36:12 si autem and they shall die without knowledge.

non audierint transibunt per gladium et consumentur in stul-

36:13 But the hypocrites in heart heap up wrath: they cry titia 36:13 simulatores et callidi prouocant iram Dei neque

not when he bindeth them.

36:14 They die in youth, and their life is among the un- clamabunt cum uincti fuerint 36:14 morietur in tempestate

clean.

36:15 He delivereth the poor in his affliction, and openeth anima eorum et uita eorum inter effeminatos 36:15 eripiet

their ears in oppression.

pauperem de angustia sua et reuelabit in tribulatione aurem

36:16 Even so would he have removed thee out of the strait eius 36:16 igitur saluabit te de ore angusto latissime et non

into a broad place, where there is no straitness; and that which should be set on thy table should be full of fatness.

habentis fundamentum subter se requies autem mensae tuae

36:17 But thou hast fulfilled the judgment of the wicked: erit plena pinguedine 36:17 causa tua quasi impii iudicata

judgment and justice take hold on thee.

36:18 Because there is wrath, beware lest he take thee est causam iudiciumque recipies 36:18 non te ergo superet

away with his stroke: then a great ransom cannot deliver thee.

ira ut aliquem opprimas nec multitudo donorum inclinet te

36:19 Will he esteem thy riches? no, not gold, nor all the 36:19 depone magnitudinem tuam absque tribulatione et om-

forces of strength.

36:20 Desire not the night, when people are cut off in their nes robustos fortitudine 36:20 ne protrahas noctem ut as-

place. 36:21 Take heed, regard not iniquity: for this hast thou cendant populi pro eis 36:21 caue ne declines ad iniquichosen rather than affliction.

36:22 Behold, God exalteth by his power: who teacheth tatem hanc enim coepisti sequi post miseriam 36:22 ecce

like him?

Deus excelsus in fortitudine sua et nullus ei similis in le-

36:23 Who hath enjoined him his way? or who can say, gislatoribus 36:23 quis poterit scrutari uias eius aut quis ei

Thou hast wrought iniquity?

36:24 Remember that thou magnify his work, which men dicere operatus es iniquitatem 36:24 memento quod igno-

behold.

36:25 Every man may see it; man may behold it afar off. res opus eius de quo cecinerunt uiri 36:25 omnes homines 36:26 Behold, God is great, and we know him not, neither uident eum unusquisque intuetur procul 36:26 ecce Deus

can the number of his years be searched out.

magnus uincens scientiam nostram numerus annorum eius

36:27 For he maketh small the drops of water: they pour inaestimabilis 36:27 qui aufert stillas pluuiae et effundit im-

down rain according to the vapour thereof: 36:28 Which the clouds do drop and distil upon man abun- bres ad instar gurgitum 36:28 qui de nubibus fluunt quae dantly.

36:29 Also can any understand the spreadings of the praetexunt cuncta desuper 36:29 si uoluerit extendere nubes

clouds, or the noise of his tabernacle? 36:30 Behold, he spreadeth his light upon it, and covereth quasi tentorium suum 36:30 et fulgurare lumine suo desuthe bottom of the sea.

36:31 For by them judgeth he the people; he giveth meat per cardines quoque maris operiet 36:31 per haec enim iu-

in abundance.

dicat populos et dat escas multis mortalibus 36:32 in mani- 36:32 With clouds he covereth the light; and commandeth

it not to shine by the cloud that cometh betwixt.

bus abscondit lucem et praecipit ei ut rursus adueniat 36:33 36:33 The noise thereof sheweth concerning it, the cattle

also concerning the vapour.

adnuntiat de ea amico suo quod possessio eius sit et ad eam possit ascendere

37:1 super hoc expauit cor meum et emotum est de loco 37:1 At this also my heart trembleth, and is moved out of

his place.

suo 37:2 audite auditionem in terrore uocis eius et sonum 37:2 Hear attentively the noise of his voice, and the sound

that goeth out of his mouth.

de ore illius procedentem 37:3 subter omnes caelos ipse 37:3 He directeth it under the whole heaven, and his light-

ning unto the ends of the earth.

considerat et lumen illius super terminos terrae 37:4 post 37:4 After it a voice roareth: he thundereth with the voice

of his excellency; and he will not stay them when his voice is heard.

eum rugiet sonitus tonabit uoce magnitudinis suae et non inuestigabitur cum audita fuerit uox eius 37:5 tonabit Deus 37:5 God thundereth marvellously with his voice; great

things doeth he, which we cannot comprehend.

in uoce sua mirabiliter qui facit magna et inscrutabilia 37:6 37:6 For he saith to the snow, Be thou on the earth; like-

wise to the small rain, and to the great rain of his strength.

qui praecipit niui ut descendat in terram et hiemis pluuiis et imbri fortitudinis suae 37:7 qui in manu omnium hominum 37:7 He sealeth up the hand of every man; that all men

may know his work.

signat ut nouerint singuli opera sua 37:8 ingredietur bes- 37:8 Then the beasts go into dens, and remain in their

places.

tia latibulum et in antro suo morabitur 37:9 ab interioribus 37:9 Out of the south cometh the whirlwind: and cold out

of the north.

egreditur tempestas et ab Arcturo frigus 37:10 flante Deo 37:10 By the breath of God frost is given: and the breadth

of the waters is straitened.

concrescit gelu et rursum latissimae funduntur aquae 37:11 37:11 Also by watering he wearieth the thick cloud: he

scattereth his bright cloud:

frumentum desiderat nubes et nubes spargunt lumen suum 37:12 quae lustrant per circuitum quocumque eas uoluntas 37:12 And it is turned round about by his counsels: that

they may do whatsoever he commandeth them upon the face of the world in the earth.

gubernantis duxerit ad omne quod praeceperit illis super fa-

ciem orbis terrarum 37:13 siue in una tribu siue in terra 37:13 He causeth it to come, whether for correction, or for

his land, or for mercy.

sua siue in quocumque loco misericordiae suae eas iusserit inueniri 37:14 ausculta haec Iob sta et considera mira- 37:14 Hearken unto this, O Job: stand still, and consider

the wondrous works of God.

cula Dei 37:15 numquid scis quando praeceperit Deus plu- 37:15 Dost thou know when God disposed them, and

caused the light of his cloud to shine?

uiis ut ostenderent lucem nubium eius 37:16 numquid nosti 37:16 Dost thou know the balancings of the clouds, the

wondrous works of him which is perfect in knowledge?

semitas nubium magnas et perfectas scientias 37:17 nonne 37:17 How thy garments are warm, when he quieteth the

earth by the south wind?

uestimenta tua calida sunt cum perflata fuerit terra austro 37:18 tu forsitan cum eo fabricatus es caelos qui solidissimi 37:18 Hast thou with him spread out the sky, which is

strong, and as a molten looking glass?

quasi aere fusi sunt 37:19 ostende nobis quid dicamus illi 37:19 Teach us what we shall say unto him; for we cannot

order our speech by reason of darkness.

nos quippe inuoluimur tenebris 37:20 quis narrabit ei quae 37:20 Shall it be told him that I speak? if a man speak,

surely he shall be swallowed up.

loquor etiam si locutus fuerit homo deuorabitur 37:21 at 37:21 And now men see not the bright light which is in the

clouds: but the wind passeth, and cleanseth them.

nunc non uident lucem subito aer cogitur in nubes et uentus transiens fugabit eas 37:22 ab aquilone aurum uenit et ad 37:22 Fair weather cometh out of the north: with God is

terrible majesty.

37:23 Touching the Almighty, we cannot find him out: he Deum formidolosa laudatio 37:23 digne eum inuenire non

is excellent in power, and in judgment, and in plenty of justice: he will not afflict.

possumus magnus fortitudine et iudicio et iustitia et enarrari

37:24 Men do therefore fear him: he respecteth not any non potest 37:24 ideo timebunt eum uiri et non audebunt

that are wise of heart.

contemplari omnes qui sibi uidentur esse sapientes

38:1 Then the LORD answered Job out of the whirlwind, 38:1 respondens autem Dominus Iob de turbine dixit 38:2

and said, 38:2 Who is this that darkeneth counsel by words without knowledge?

quis est iste inuoluens sententias sermonibus inperitis 38:3

38:3 Gird up now thy loins like a man; for I will demand

of thee, and answer thou me. accinge sicut uir lumbos tuos interrogabo te et responde 38:4 Where wast thou when I laid the foundations of the mihi 38:4 ubi eras quando ponebam fundamenta terrae in-

earth? declare, if thou hast understanding.

38:5 Who hath laid the measures thereof, if thou knowest? dica mihi si habes intellegentiam 38:5 quis posuit mensuras

or who hath stretched the line upon it?

38:6 Whereupon are the foundations thereof fastened? or eius si nosti uel quis tetendit super eam lineam 38:6 super

who laid the corner stone thereof;

quo bases illius solidatae sunt aut quis dimisit lapidem an-

38:7 When the morning stars sang together, and all the gularem eius 38:7 cum me laudarent simul astra matutina

sons of God shouted for joy?

38:8 Or who shut up the sea with doors, when it brake et iubilarent omnes filii Dei 38:8 quis conclusit ostiis mare

forth, as if it had issued out of the womb? 38:9 When I made the cloud the garment thereof, and thick quando erumpebat quasi de uulua procedens 38:9 cum podarkness a swaddlingband for it,

nerem nubem uestimentum eius et caligine illud quasi pan-

38:10 And brake up for it my decreed place, and set bars nis infantiae obuoluerem 38:10 circumdedi illud terminis

and doors,

38:11 And said, Hitherto shalt thou come, but no further: meis et posui uectem et ostia 38:11 et dixi usque huc uenies

and here shall thy proud waves be stayed?

et non procedes amplius et hic confringes tumentes fluctus

38:12 Hast thou commanded the morning since thy days; tuos 38:12 numquid post ortum tuum praecepisti diluculo et

and caused the dayspring to know his place;

38:13 That it might take hold of the ends of the earth, that ostendisti aurorae locum suum 38:13 et tenuisti concutiens

the wicked might be shaken out of it?

38:14 It is turned as clay to the seal; and they stand as a extrema terrae et excussisti impios ex ea 38:14 restituetur

garment.

38:15 And from the wicked their light is withholden, and ut lutum signaculum et stabit sicut uestimentum 38:15 au-

the high arm shall be broken.

feretur ab impiis lux sua et brachium excelsum confringetur

38:16 Hast thou entered into the springs of the sea? or hast 38:16 numquid ingressus es profunda maris et in nouissimis

thou walked in the search of the depth?

38:17 Have the gates of death been opened unto thee? or abyssis deambulasti 38:17 numquid apertae tibi sunt por-

hast thou seen the doors of the shadow of death?

38:18 Hast thou perceived the breadth of the earth? de- tae mortis et ostia tenebrosa uidisti 38:18 numquid consi-

clare if thou knowest it all.

38:19 Where is the way where light dwelleth? and as for derasti latitudines terrae indica mihi si nosti omnia 38:19

darkness, where is the place thereof, 38:20 That thou shouldest take it to the bound thereof, and in qua uia habitet lux et tenebrarum quis locus sit 38:20 that thou shouldest know the paths to the house thereof?

ut ducas unumquodque ad terminos suos et intellegas se-

38:21 Knowest thou it, because thou wast then born? or mitas domus eius 38:21 sciebas tunc quod nasciturus es-

because the number of thy days is great? 38:22 Hast thou entered into the treasures of the snow? or ses et numerum dierum tuorum noueras 38:22 numquid inhast thou seen the treasures of the hail,

gressus es thesauros niuis aut thesauros grandinis aspexisti

38:23 Which I have reserved against the time of trouble, 38:23 quae praeparaui in tempus hostis in diem pugnae et against the day of battle and war? belli 38:24 per quam uiam spargitur lux diuiditur aestus 38:24 By what way is the light parted, which scattereth the

east wind upon the earth?

super terram 38:25 quis dedit uehementissimo imbri cur- 38:25 Who hath divided a watercourse for the overflowing

of waters, or a way for the lightning of thunder;

sum et uiam sonantis tonitrui 38:26 ut plueret super terram 38:26 To cause it to rain on the earth, where no man is; on

the wilderness, wherein there is no man;

absque homine in deserto ubi nullus mortalium commoratur 38:27 ut impleret inuiam et desolatam et produceret herbas 38:27 To satisfy the desolate and waste ground; and to

cause the bud of the tender herb to spring forth?

uirentes 38:28 quis est pluuiae pater uel quis genuit stil- 38:28 Hath the rain a father? or who hath begotten the

drops of dew?

las roris 38:29 de cuius utero egressa est glacies et gelu 38:29 Out of whose womb came the ice? and the hoary

frost of heaven, who hath gendered it?

de caelo quis genuit 38:30 in similitudinem lapidis aquae 38:30 The waters are hid as with a stone, and the face of

the deep is frozen.

durantur et superficies abyssi constringitur 38:31 numquid 38:31 Canst thou bind the sweet influences of Pleiades, or

loose the bands of Orion?

coniungere ualebis micantes stellas Pliadis aut gyrum Arcturi poteris dissipare 38:32 numquid producis luciferum in 38:32 Canst thou bring forth Mazzaroth in his season? or

canst thou guide Arcturus with his sons?

tempore suo et uesperum super filios terrae consurgere facis 38:33 numquid nosti ordinem caeli et pones rationem eius in 38:33 Knowest thou the ordinances of heaven? canst thou

set the dominion thereof in the earth?

terra 38:34 numquid eleuabis in nebula uocem tuam et im- 38:34 Canst thou lift up thy voice to the clouds, that abun-

dance of waters may cover thee?

petus aquarum operiet te 38:35 numquid mittes fulgura et 38:35 Canst thou send lightnings, that they may go and

say unto thee, Here we are?

ibunt et reuertentia dicent tibi adsumus 38:36 quis posuit 38:36 Who hath put wisdom in the inward parts? or who

hath given understanding to the heart?

in uisceribus hominis sapientiam uel quis dedit gallo intellegentiam 38:37 quis enarrauit caelorum rationem et con- 38:37 Who can number the clouds in wisdom? or who can

stay the bottles of heaven,

centum caeli quis dormire faciet 38:38 quando fundebatur 38:38 When the dust groweth into hardness, and the clods

cleave fast together?

puluis in terram et glebae conpingebantur 38:39 numquid 38:39 Wilt thou hunt the prey for the lion? or fill the ap-

petite of the young lions,

capies leaenae praedam et animam catulorum eius implebis 38:40 quando cubant in antris et in specubus insidiantur 38:40 When they couch in their dens, and abide in the

covert to lie in wait?

38:41 quis praeparat coruo escam suam quando pulli eius ad 38:41 Who provideth for the raven his food? when his

young ones cry unto God, they wander for lack of meat.

Deum clamant uagantes eo quod non habeant cibos

39:1 numquid nosti tempus partus hibicum in petris 39:1 Knowest thou the time when the wild goats of the

rock bring forth? or canst thou mark when the hinds do calve?

uel parturientes ceruas obseruasti 39:2 dinumerasti menses

39:2 Canst thou number the months that they fulfil? or knowest thou the time when they bring forth?

conceptus earum et scisti tempus partus earum 39:3 incur- 39:3 They bow themselves, they bring forth their young

ones, they cast out their sorrows.

uantur ad fetum et pariunt et rugitus emittunt 39:4 sepa- 39:4 Their young ones are in good liking, they grow up

with corn; they go forth, and return not unto them.

rantur filii earum pergunt ad pastum egrediuntur et non reuertuntur ad eas 39:5 quis dimisit onagrum liberum et uin- 39:5 Who hath sent out the wild ass free? or who hath

loosed the bands of the wild ass?

cula eius quis soluit 39:6 cui dedi in solitudine domum et 39:6 Whose house I have made the wilderness, and the

barren land his dwellings.

tabernacula eius in terra salsuginis 39:7 contemnit multi- 39:7 He scorneth the multitude of the city, neither re-

gardeth he the crying of the driver.

tudinem ciuitatis clamorem exactoris non audit 39:8 cir- 39:8 The range of the mountains is his pasture, and he

searcheth after every green thing.

cumspicit montes pascuae suae et uirentia quaeque perqui-

39:9 Will the unicorn be willing to serve thee, or abide by rit 39:9 numquid uolet rinoceros seruire tibi aut morabi-

thy crib?

39:10 Canst thou bind the unicorn with his band in the tur ad praesepe tuum 39:10 numquid alligabis rinocerota

furrow? or will he harrow the valleys after thee?

ad arandum loro tuo aut confringet glebas uallium post

39:11 Wilt thou trust him, because his strength is great? or te 39:11 numquid fiduciam habebis in magna fortitudine wilt thou leave thy labour to him?

39:12 Wilt thou believe him, that he will bring home thy eius et derelinques ei labores tuos 39:12 numquid credes

seed, and gather it into thy barn?

ei quoniam reddat sementem tibi et aream tuam congreget

39:13 Gavest thou the goodly wings unto the peacocks? or 39:13 pinna strutionum similis est pinnis herodii et accipitris

wings and feathers unto the ostrich?

39:14 Which leaveth her eggs in the earth, and warmeth 39:14 quando derelinquit in terra oua sua tu forsitan in pul-

them in dust,

39:15 And forgetteth that the foot may crush them, or that uere calefacis ea 39:15 obliuiscitur quod pes conculcet ea

the wild beast may break them.

39:16 She is hardened against her young ones, as though aut bestiae agri conterant 39:16 duratur ad filios suos quasi

they were not her’s: her labour is in vain without fear;

39:17 Because God hath deprived her of wisdom, neither non sint sui frustra laborauit nullo timore cogente 39:17 pri-

hath he imparted to her understanding.

uauit enim eam Deus sapientia nec dedit illi intellegentiam

39:18 What time she lifteth up herself on high, she scor- 39:18 cum tempus fuerit in altum alas erigit deridet equi-

neth the horse and his rider.

39:19 Hast thou given the horse strength? hast thou tem et ascensorem eius 39:19 numquid praebebis equo for-

clothed his neck with thunder?

39:20 Canst thou make him afraid as a grasshopper? the titudinem aut circumdabis collo eius hinnitum 39:20 num-

glory of his nostrils is terrible.

quid suscitabis eum quasi lucustas gloria narium eius ter-

39:21 He paweth in the valley, and rejoiceth in his ror 39:21 terram ungula fodit exultat audacter in occursum

strength: he goeth on to meet the armed men.

39:22 He mocketh at fear, and is not affrighted; neither pergit armatis 39:22 contemnit pauorem nec cedit gladio

turneth he back from the sword. 39:23 The quiver rattleth against him, the glittering spear 39:23 super ipsum sonabit faretra uibrabit hasta et clypeus and the shield.

39:24 He swalloweth the ground with fierceness and rage: 39:24 feruens et fremens sorbet terram nec reputat tubae so-

neither believeth he that it is the sound of the trumpet.

39:25 He saith among the trumpets, Ha, ha; and he nare clangorem 39:25 ubi audierit bucinam dicet ua procul

smelleth the battle afar off, the thunder of the captains, and the shouting.

odoratur bellum exhortationem ducum et ululatum exerci-

39:26 Doth the hawk fly by thy wisdom, and stretch her tus 39:26 numquid per sapientiam tuam plumescit accipi-

wings toward the south?

39:27 Doth the eagle mount up at thy command, and make ter expandens alas suas ad austrum 39:27 aut ad praecep-

her nest on high?

tum tuum eleuabitur aquila et in arduis ponet nidum suum

39:28 She dwelleth and abideth on the rock, upon the crag 39:28 in petris manet et in praeruptis silicibus commoratur at-

of the rock, and the strong place.

39:29 From thence she seeketh the prey, and her eyes be- que inaccessis rupibus 39:29 inde contemplatur escam et de

hold afar off.

39:30 Her young ones also suck up blood: and where the longe oculi eius prospiciunt 39:30 pulli eius lambent san-

slain are, there is she.

39:31 Moreover the LORD answered Job, and said, guinem et ubicumque cadauer fuerit statim adest 39:31 et

39:32 Shall he that contendeth with the Almighty instruct adiecit Dominus et locutus est ad Iob 39:32 numquid qui him? he that reproveth God, let him answer it.

contendit cum Deo tam facile conquiescit utique qui arguit

39:33 Then Job answered the LORD, and said, Deum debet respondere ei 39:33 respondens autem Iob Domino dixit 39:34 qui leuiter locutus sum respondere quid 39:34 Behold, I am vile; what shall I answer thee? I will

lay mine hand upon my mouth.

possum manum meam ponam super os meum 39:35 unum 39:35 Once have I spoken; but I will not answer: yea,

twice; but I will proceed no further.

locutus sum quod utinam non dixissem et alterum quibus ultra non addam

40:1 respondens autem Dominus Iob de turbine ait 40:2 40:1 Then answered the LORD unto Job out of the whirl-

wind, and said,

40:2 Gird up thy loins now like a man: I will demand of thee, and declare thou unto me.

accinge sicut uir lumbos tuos interrogabo te et indica mihi 40:3 numquid irritum facies iudicium meum et condemnabis 40:3 Wilt thou also disannul my judgment? wilt thou con-

demn me, that thou mayest be righteous?

me ut tu iustificeris 40:4 et si habes brachium sicut Deus 40:4 Hast thou an arm like God? or canst thou thunder

with a voice like him?

et si uoce simili tonas 40:5 circumda tibi decorem et in su- 40:5 Deck thyself now with majesty and excellency; and

array thyself with glory and beauty.

blime erigere et esto gloriosus et speciosis induere uestibus 40:6 disperge superbos furore tuo et respiciens omnem arro- 40:6 Cast abroad the rage of thy wrath: and behold every

one that is proud, and abase him.

gantem humilia 40:7 respice cunctos superbos et confunde 40:7 Look on every one that is proud, and bring him low;

and tread down the wicked in their place.

eos et contere impios in loco suo 40:8 absconde eos in pul- 40:8 Hide them in the dust together; and bind their faces

in secret.

uere simul et facies eorum demerge in foueam 40:9 et ego 40:9 Then will I also confess unto thee that thine own right

hand can save thee.

confitebor quod saluare te possit dextera tua 40:10 ecce Be- 40:10 Behold now behemoth, which I made with thee; he

eateth grass as an ox.

hemoth quem feci tecum faenum quasi bos comedet 40:11 40:11 Lo now, his strength is in his loins, and his force is

in the navel of his belly.

fortitudo eius in lumbis eius et uirtus illius in umbilicis uentris eius 40:12 constringit caudam suam quasi cedrum nerui 40:12 He moveth his tail like a cedar: the sinews of his

stones are wrapped together.

testiculorum eius perplexi sunt 40:13 ossa eius uelut fistu- 40:13 His bones are as strong pieces of brass; his bones

are like bars of iron.

lae aeris cartilago illius quasi lamminae ferreae 40:14 ipse 40:14 He is the chief of the ways of God: he that made

him can make his sword to approach unto him.

principium est uiarum Dei qui fecit eum adplicabit gladium eius 40:15 huic montes herbas ferunt omnes bestiae agri lu- 40:15 Surely the mountains bring him forth food, where

all the beasts of the field play.

dent ibi 40:16 sub umbra dormit in secreto calami et locis 40:16 He lieth under the shady trees, in the covert of the

reed, and fens.

humentibus 40:17 protegunt umbrae umbram eius circum- 40:17 The shady trees cover him with their shadow; the

willows of the brook compass him about.

dabunt eum salices torrentis 40:18 ecce absorbebit fluuium 40:18 Behold, he drinketh up a river, and hasteth not: he

trusteth that he can draw up Jordan into his mouth.

et non mirabitur habet fiduciam quod influat Iordanis in os

eius 40:19 in oculis eius quasi hamo capiet eum et in sudi- 40:19 He taketh it with his eyes: his nose pierceth through

snares.

bus perforabit nares eius 40:20 an extrahere poteris Leuia- 40:20 Canst thou draw out leviathan with an hook? or his

tongue with a cord which thou lettest down?

than hamo et fune ligabis linguam eius 40:21 numquid po- 40:21 Canst thou put an hook into his nose? or bore his

jaw through with a thorn?

nes circulum in naribus eius et armilla perforabis maxillam eius 40:22 numquid multiplicabit ad te preces aut loque- 40:22 Will he make many supplications unto thee? will he

speak soft words unto thee?

tur tibi mollia 40:23 numquid feriet tecum pactum et ac- 40:23 Will he make a covenant with thee? wilt thou take

him for a servant for ever?

cipies eum seruum sempiternum 40:24 numquid inludes ei 40:24 Wilt thou play with him as with a bird? or wilt thou

bind him for thy maidens?

40:25 Shall the companions make a banquet of him? shall quasi aui aut ligabis illum ancillis tuis 40:25 concident eum

they part him among the merchants?

40:26 Canst thou fill his skin with barbed irons? or his amici diuident illum negotiatores 40:26 numquid implebis

head with fish spears?

40:27 Lay thine hand upon him, remember the battle, do sagenas pelle eius et gurgustium piscium capite illius 40:27

no more.

pone super eum manum tuam memento belli nec ultra ad-

40:28 Behold, the hope of him is in vain: shall not one be das loqui 40:28 ecce spes eius frustrabitur eum et uidenti-

cast down even at the sight of him?

bus cunctis praecipitabitur

41:1 None is so fierce that dare stir him up: who then is 41:1 non quasi crudelis suscitabo eum quis enim resistere

able to stand before me?

41:2 Who hath prevented me, that I should repay him? potest uultui meo 41:2 quis ante dedit mihi ut reddam ei

whatsoever is under the whole heaven is mine. 41:3 I will not conceal his parts, nor his power, nor his omnia quae sub caelo sunt mea sunt 41:3 non parcam ei et comely proportion.

41:4 Who can discover the face of his garment? or who uerbis potentibus et ad deprecandum conpositis 41:4 quis

can come to him with his double bridle?

reuelauit faciem indumenti eius et in medium oris eius quis

41:5 Who can open the doors of his face? his teeth are intrabit 41:5 portas uultus eius quis aperiet per gyrum de-

terrible round about.

41:6 His scales are his pride, shut up together as with a ntium eius formido 41:6 corpus illius quasi scuta fusilia et

close seal.

41:7 One is so near to another, that no air can come be- conpactum squamis se prementibus 41:7 una uni coniun-

tween them. 41:8 They are joined one to another, they stick together, gitur et ne spiraculum quidem incedit per eas 41:8 una althat they cannot be sundered.

41:9 By his neesings a light doth shine, and his eyes are teri adherebunt et tenentes se nequaquam separabuntur 41:9

like the eyelids of the morning.

sternutatio eius splendor ignis et oculi eius ut palpebrae di-

41:10 Out of his mouth go burning lamps, and sparks of luculi 41:10 de ore eius lampades procedunt sicut taedae

fire leap out. 41:11 Out of his nostrils goeth smoke, as out of a seething ignis accensae 41:11 de naribus eius procedit fumus sicpot or caldron.

41:12 His breath kindleth coals, and a flame goeth out of ut ollae succensae atque feruentis 41:12 halitus eius pru-

his mouth. 41:13 In his neck remaineth strength, and sorrow is turned nas ardere facit et flamma de ore eius egreditur 41:13 in into joy before him.

collo eius morabitur fortitudo et faciem eius praecedet ege-

41:14 The flakes of his flesh are joined together: they are stas 41:14 membra carnium eius coherentia sibi mittet con-

firm in themselves; they cannot be moved. 41:15 His heart is as firm as a stone; yea, as hard as a piece tra eum fulmina et ad locum alium non ferentur 41:15 cor of the nether millstone.

eius indurabitur quasi lapis et stringetur quasi malleatoris in-

41:16 When he raiseth up himself, the mighty are afraid: cus 41:16 cum sublatus fuerit timebunt angeli et territi pur-

by reason of breakings they purify themselves.

41:17 The sword of him that layeth at him cannot hold: gabuntur 41:17 cum adprehenderit eum gladius subsistere

the spear, the dart, nor the habergeon.

41:18 He esteemeth iron as straw, and brass as rotten non poterit neque hasta neque torax 41:18 reputabit enim

wood.

41:19 The arrow cannot make him flee: slingstones are quasi paleas ferrum et quasi lignum putridum aes 41:19 non

turned with him into stubble.

fugabit eum uir sagittarius in stipulam uersi sunt ei lapides

41:20 Darts are counted as stubble: he laugheth at the fundae 41:20 quasi stipulam aestimabit malleum et deride-

shaking of a spear.

41:21 Sharp stones are under him: he spreadeth sharp bit uibrantem hastam 41:21 sub ipso erunt radii solis sternet

pointed things upon the mire.

41:22 He maketh the deep to boil like a pot: he maketh the sibi aurum quasi lutum 41:22 feruescere faciet quasi ollam

sea like a pot of ointment.

profundum mare ponet quasi cum unguenta bulliunt 41:23 41:23 He maketh a path to shine after him; one would

think the deep to be hoary.

post eum lucebit semita aestimabit abyssum quasi senescentem 41:24 non est super terram potestas quae conparetur ei 41:24 Upon earth there is not his like, who is made without

fear.

qui factus est ut nullum timeret 41:25 omne sublime uidet 41:25 He beholdeth all high things: he is a king over all

the children of pride.

ipse est rex super uniuersos filios superbiae

42:1 respondens autem Iob Domino dixit 42:2 scio quia 42:1 Then Job answered the LORD, and said,

42:2 I know that thou canst do every thing, and that no thought can be withholden from thee.

omnia potes et nulla te latet cogitatio 42:3 quis est iste qui

42:3 Who is he that hideth counsel without knowledge? therefore have I uttered that I understood not; things too

celat consilium absque scientia ideo insipienter locutus sum _{wonderful for me, which I knew not.} et quae ultra modum excederent scientiam meam 42:4 audi 42:4 Hear, I beseech thee, and I will speak: I will demand

of thee, and declare thou unto me.

et ego loquar interrogabo et ostende mihi 42:5 auditu au- 42:5 I have heard of thee by the hearing of the ear: but

now mine eye seeth thee.

ris audiui te nunc autem oculus meus uidet te 42:6 idcirco 42:6 Wherefore I abhor myself, and repent in dust and

ashes.

ipse me reprehendo et ago paenitentiam in fauilla et cinere 42:7 postquam autem locutus est Dominus uerba haec ad Iob 42:7 And it was so, that after the LORD had spoken these

words unto Job, the LORD said to Eliphaz the Temanite, My wrath is kindled against thee, and against thy two friends: for ye have not spoken of me the thing that is right, as my servant Job hath.

dixit ad Eliphaz Themaniten iratus est furor meus in te et in duos amicos tuos quoniam non estis locuti coram me rectum sicut seruus meus Iob 42:8 sumite igitur uobis septem 42:8 Therefore take unto you now seven bullocks and

seven rams, and go to my servant Job, and offer up for yourselves a burnt offering; and my servant Job shall pray for you: for him will I accept: lest I deal with you after your folly, in that ye have not spoken of me the thing which is right, like my servant Job.

tauros et septem arietes et ite ad seruum meum Iob et offerte holocaustum pro uobis Iob autem seruus meus orabit pro uobis faciem eius suscipiam ut non uobis inputetur stultitia neque enim locuti estis ad me recta sicut seruus meus Iob 42:9 abierunt ergo Eliphaz Themanites et Baldad Suites 42:9 So Eliphaz the Temanite and Bildad the Shuhite and

Zophar the Naamathite went, and did according as the LORD commanded them: the LORD also accepted Job.

et Sophar Naamathites et fecerunt sicut locutus fuerat ad eos Dominus et suscepit Dominus faciem Iob 42:10 Dominus 42:10 And the LORD turned the captivity of Job, when he

prayed for his friends: also the LORD gave Job twice as much as he had before.

quoque conuersus est ad paenitentiam Iob cum oraret ille pro amicis suis et addidit Dominus omnia quaecumque fuerant Iob duplicia 42:11 uenerunt autem ad eum omnes fratres sui 42:11 Then came there unto him all his brethren, and all

his sisters, and all they that had been of his acquaintance before, and did eat bread with him in his house: and they bemoaned him, and comforted him over all the evil that the LORD had brought upon him: every man also gave him a piece of money, and every one an earring of gold.

et uniuersae sorores suae et cuncti qui nouerant eum prius et comederunt cum eo panem in domo eius et mouerunt super eum caput et consolati sunt eum super omni malo quod intulerat Dominus super eum et dederunt ei unusquisque ouem unam et inaurem auream unam 42:12 Dominus autem be- 42:12 So the LORD blessed the latter end of Job more

than his beginning: for he had fourteen thousand sheep, and six thousand camels, and a thousand yoke of oxen,

nedixit nouissimis Iob magis quam principio eius et facta and a thousand she asses. sunt ei quattuordecim milia ouium et sex milia camelorum

Biblia Sacra Vulgata

42:13 He had also seven sons and three daughters. et mille iuga boum et mille asinae 42:13 et fuerunt ei sep-42:14 And he called the name of the first, Jemima; and tem filii et filiae tres 42:14 et uocauit nomen unius Diem et

the name of the second, Kezia; and the name of the third, Kerenhappuch.

nomen secundae Cassia et nomen tertiae Cornu stibii 42:15

42:15 And in all the land were no women found so fair as the daughters of Job: and their father gave them inheritance among their brethren.

non sunt autem inuentae mulieres speciosae sicut filiae Iob in uniuersa terra deditque eis pater suus hereditatem inter

42:16 After this lived Job an hundred and forty years, and fratres earum 42:16 uixit autem Iob post haec centum quasaw his sons, and his sons’ sons, even four generations.

Liber Prouerbiorum

draginta annis et uidit filios suos et filios filiorum suorum usque ad quartam generationem et mortuus est senex et plenus dierum 1:1 parabolae Salomonis filii Dauid regis Israhel 1:2 ad 1:1 The proverbs of Solomon the son of David, king of

Israel;

1:2 To know wisdom and instruction; to perceive the words of understanding;

sciendam sapientiam et disciplinam 1:3 ad intellegenda

1:3 To receive the instruction of wisdom, justice, and judg-

uerba prudentiae et suscipiendam eruditionem doctrinae iu- _{ment, and equity;} stitiam et iudicium et aequitatem 1:4 ut detur paruulis astu- 1:4 To give subtilty to the simple, to the young man knowl-

edge and discretion.

tia adulescenti scientia et intellectus 1:5 audiens sapiens sa- 1:5 A wise man will hear, and will increase learning; and

a man of understanding shall attain unto wise counsels:

pientior erit et intellegens gubernacula possidebit 1:6 ani- 1:6 To understand a proverb, and the interpretation; the

words of the wise, and their dark sayings.

maduertet parabolam et interpretationem uerba sapientium et enigmata eorum 1:7 timor Domini principium scientiae 1:7 The fear of the LORD is the beginning of knowledge:

but fools despise wisdom and instruction.

sapientiam atque doctrinam stulti despiciunt 1:8 audi fili 1:8 My son, hear the instruction of thy father, and forsake

not the law of thy mother:

mi disciplinam patris tui et ne dimittas legem matris tuae 1:9 ut addatur gratia capiti tuo et torques collo tuo 1:10 fili 1:9 For they shall be an ornament of grace unto thy head,

and chains about thy neck.

1:10 My son, if sinners entice thee, consent thou not.

mi si te lactauerint peccatores ne adquiescas 1:11 si dixe-

1:11 If they say, Come with us, let us lay wait for blood, let us lurk privily for the innocent without cause:

rint ueni nobiscum insidiemur sanguini abscondamus tendiculas contra insontem frustra 1:12 degluttiamus eum sic- 1:12 Let us swallow them up alive as the grave; and whole,

as those that go down into the pit:

ut infernus uiuentem et integrum quasi descendentem in la-

cum 1:13 omnem pretiosam substantiam repperiemus imp- 1:13 We shall find all precious substance, we shall fill our

houses with spoil:

lebimus domos nostras spoliis 1:14 sortem mitte nobiscum 1:14 Cast in thy lot among us; let us all have one purse:

marsuppium unum sit omnium nostrum 1:15 fili mi ne am- 1:15 My son, walk not thou in the way with them; refrain

thy foot from their path:

bules cum eis prohibe pedem tuum a semitis eorum 1:16 1:16 For their feet run to evil, and make haste to shed

blood.

pedes enim illorum ad malum currunt et festinant ut effundant sanguinem 1:17 frustra autem iacitur rete ante oculos 1:17 Surely in vain the net is spread in the sight of any

bird.

pinnatorum 1:18 ipsique contra sanguinem suum insidian- 1:18 And they lay wait for their own blood; they lurk priv-

ily for their own lives.

tur et moliuntur fraudes contra animas suas 1:19 sic semi- 1:19 So are the ways of every one that is greedy of gain;

which taketh away the life of the owners thereof.

tae omnis auari animas possidentium rapiunt 1:20 sapientia 1:20 Wisdom crieth without; she uttereth her voice in the

streets:

foris praedicat in plateis dat uocem suam 1:21 in capite tur- 1:21 She crieth in the chief place of concourse, in the

openings of the gates: in the city she uttereth her words, saying,

barum clamitat in foribus portarum urbis profert uerba sua dicens 1:22 usquequo paruuli diligitis infantiam et stulti ea 1:22 How long, ye simple ones, will ye love simplicity?

and the scorners delight in their scorning, and fools hate knowledge?

quae sibi sunt noxia cupiunt et inprudentes odibunt scientiam 1:23 conuertimini ad correptionem meam en proferam 1:23 Turn you at my reproof: behold, I will pour out my

spirit unto you, I will make known my words unto you.

uobis spiritum meum et ostendam uerba mea 1:24 quia uo- 1:24 Because I have called, and ye refused; I have

stretched out my hand, and no man regarded;

caui et rennuistis extendi manum meam et non fuit qui aspiceret 1:25 despexistis omne consilium meum et increpa- 1:25 But ye have set at nought all my counsel, and would

none of my reproof:

tiones meas neglexistis 1:26 ego quoque in interitu uestro 1:26 I also will laugh at your calamity; I will mock when

your fear cometh;

ridebo et subsannabo cum uobis quod timebatis aduenerit

1:27 When your fear cometh as desolation, and your de- 1:27 cum inruerit repentina calamitas et interitus quasi tem-

struction cometh as a whirlwind; when distress and anguish cometh upon you.

pestas ingruerit quando uenerit super uos tribulatio et angu-

1:28 Then shall they call upon me, but I will not answer; stia 1:28 tunc inuocabunt me et non exaudiam mane conthey shall seek me early, but they shall not find me: 1:29 For that they hated knowledge, and did not choose surgent et non inuenient me 1:29 eo quod exosam habuethe fear of the LORD:

1:30 They would none of my counsel: they despised all rint disciplinam et timorem Domini non susceperint 1:30

my reproof.

nec adquieuerint consilio meo et detraxerint uniuersae cor-

1:31 Therefore shall they eat of the fruit of their own way, reptioni meae 1:31 comedent igitur fructus uiae suae suis-

and be filled with their own devices.

1:32 For the turning away of the simple shall slay them, que consiliis saturabuntur 1:32 auersio paruulorum interfi-

and the prosperity of fools shall destroy them.

1:33 But whoso hearkeneth unto me shall dwell safely, and ciet eos et prosperitas stultorum perdet illos 1:33 qui autem

shall be quiet from fear of evil.

me audierit absque terrore requiescet et abundantia perfruetur malorum timore sublato

2:1 My son, if thou wilt receive my words, and hide my 2:1 fili mi si susceperis sermones meos et mandata mea ab-

commandments with thee;

2:2 So that thou incline thine ear unto wisdom, and apply sconderis penes te 2:2 ut audiat sapientiam auris tua inclina

thine heart to understanding; 2:3 Yea, if thou criest after knowledge, and liftest up thy cor tuum ad noscendam prudentiam 2:3 si enim sapientiam voice for understanding;

2:4 If thou seekest her as silver, and searchest for her as inuocaueris et inclinaueris cor tuum prudentiae 2:4 si quae-

for hid treasures;

sieris eam quasi pecuniam et sicut thesauros effoderis illam

2:5 Then shalt thou understand the fear of the LORD, and 2:5 tunc intelleges timorem Domini et scientiam Dei inue-

find the knowledge of God.

2:6 For the LORD giveth wisdom: out of his mouth nies 2:6 quia Dominus dat sapientiam et ex ore eius scien-

cometh knowledge and understanding.

2:7 He layeth up sound wisdom for the righteous: he is a tia et prudentia 2:7 custodiet rectorum salutem et proteget

buckler to them that walk uprightly.

2:8 He keepeth the paths of judgment, and preserveth the gradientes simpliciter 2:8 seruans semitas iustitiae et uias

way of his saints.

2:9 Then shalt thou understand righteousness, and judg- sanctorum custodiens 2:9 tunc intelleges iustitiam et iudi-

ment, and equity; yea, every good path. 2:10 When wisdom entereth into thine heart, and knowl- cium et aequitatem et omnem semitam bonam 2:10 si inedge is pleasant unto thy soul;

trauerit sapientia cor tuum et scientia animae tuae placue-

2:11 Discretion shall preserve thee, understanding shall rit 2:11 consilium custodiet te prudentia seruabit te 2:12 keep thee: 2:12 To deliver thee from the way of the evil man, from the man that speaketh froward things;

ut eruaris de uia mala ab homine qui peruersa loquitur 2:13

2:13 Who leave the paths of uprightness, to walk in the

ways of darkness; qui relinquunt iter rectum et ambulant per uias tenebrosas 2:14 Who rejoice to do evil, and delight in the frowardness 2:14 qui laetantur cum malefecerint et exultant in rebus pes-

of the wicked;

2:15 Whose ways are crooked, and they froward in their simis 2:15 quorum uiae peruersae et infames gressus eorum

paths:

2:16 To deliver thee from the strange woman, even from 2:16 ut eruaris a muliere aliena et ab extranea quae mollit ser-

the stranger which flattereth with her words; 2:17 Which forsaketh the guide of her youth, and forget- mones suos 2:17 et relinquit ducem pubertatis suae 2:18 teth the covenant of her God. 2:18 For her house inclineth unto death, and her paths unto the dead.

et pacti Dei sui oblita est inclinata est enim ad mortem do-

2:19 None that go unto her return again, neither take they mus eius et ad impios semitae ipsius 2:19 omnes qui ingrehold of the paths of life.

diuntur ad eam non reuertentur nec adprehendent semitas uitae 2:20 ut ambules in uia bona et calles iustorum custo- 2:20 That thou mayest walk in the way of good men, and

keep the paths of the righteous.

dias 2:21 qui enim recti sunt habitabunt in terra et simplices 2:21 For the upright shall dwell in the land, and the perfect

shall remain in it.

permanebunt in ea 2:22 impii uero de terra perdentur et qui 2:22 But the wicked shall be cut off from the earth, and

the transgressors shall be rooted out of it.

inique agunt auferentur ex ea

3:1 fili mi ne obliuiscaris legis meae et praecepta mea 3:1 My son, forget not my law; but let thine heart keep my

commandments:

custodiat cor tuum 3:2 longitudinem enim dierum et an- 3:2 For length of days, and long life, and peace, shall they

add to thee.

nos uitae et pacem adponent tibi 3:3 misericordia et ueritas 3:3 Let not mercy and truth forsake thee: bind them about

thy neck; write them upon the table of thine heart:

non te deserant circumda eas gutturi tuo et describe in tabulis cordis tui 3:4 et inuenies gratiam et disciplinam bonam 3:4 So shalt thou find favour and good understanding in

the sight of God and man.

coram Deo et hominibus 3:5 habe fiduciam in Domino ex 3:5 Trust in the LORD with all thine heart; and lean not

unto thine own understanding.

toto corde tuo et ne innitaris prudentiae tuae 3:6 in omni- 3:6 In all thy ways acknowledge him, and he shall direct

thy paths.

bus uiis tuis cogita illum et ipse diriget gressus tuos 3:7 3:7 Be not wise in thine own eyes: fear the LORD, and

depart from evil.

ne sis sapiens apud temet ipsum time Dominum et recede a malo 3:8 sanitas quippe erit umbilico tuo et inrigatio os- 3:8 It shall be health to thy navel, and marrow to thy bones. suum tuorum 3:9 honora Dominum de tua substantia et de 3:9 Honour the LORD with thy substance, and with the

firstfruits of all thine increase:

primitiis omnium frugum tuarum 3:10 et implebuntur hor- 3:10 So shall thy barns be filled with plenty, and thy

presses shall burst out with new wine.

rea tua saturitate et uino torcularia redundabunt 3:11 dis- 3:11 My son, despise not the chastening of the LORD;

neither be weary of his correction:

ciplinam Domini fili mi ne abicias nec deficias cum ab eo corriperis 3:12 quem enim diligit Dominus corripit et quasi 3:12 For whom the LORD loveth he correcteth; even as a

father the son in whom he delighteth.

pater in filio conplacet sibi 3:13 beatus homo qui inuenit 3:13 Happy is the man that findeth wisdom, and the man

that getteth understanding.

sapientiam et qui affluit prudentia 3:14 melior est adquisi- 3:14 For the merchandise of it is better than the merchan-

dise of silver, and the gain thereof than fine gold.

tio eius negotiatione argenti et auro primo fructus eius 3:15 3:15 She is more precious than rubies: and all the things

thou canst desire are not to be compared unto her.

pretiosior est cunctis opibus et omnia quae desiderantur huic non ualent conparari 3:16 longitudo dierum in dextera eius 3:16 Length of days is in her right hand; and in her left

hand riches and honour.

in sinistra illius diuitiae et gloria 3:17 uiae eius uiae pulch- 3:17 Her ways are ways of pleasantness, and all her paths

are peace.

rae et omnes semitae illius pacificae 3:18 lignum uitae est 3:18 She is a tree of life to them that lay hold upon her:

and happy is every one that retaineth her.

his qui adprehenderint eam et qui tenuerit eam beatus 3:19 3:19 The LORD by wisdom hath founded the earth; by

understanding hath he established the heavens.

Dominus sapientia fundauit terram stabiliuit caelos prudentia 3:20 sapientia illius eruperunt abyssi et nubes rore con- 3:20 By his knowledge the depths are broken up, and the

clouds drop down the dew.

crescunt 3:21 fili mi ne effluant haec ab oculis tuis custodi 3:21 My son, let not them depart from thine eyes: keep

sound wisdom and discretion:

legem atque consilium 3:22 et erit uita animae tuae et gra- 3:22 So shall they be life unto thy soul, and grace to thy

neck.

tia faucibus tuis 3:23 tunc ambulabis fiducialiter in uia tua 3:23 Then shalt thou walk in thy way safely, and thy foot

shall not stumble.

3:24 When thou liest down, thou shalt not be afraid: yea, et pes tuus non inpinget 3:24 si dormieris non timebis quie-

thou shalt lie down, and thy sleep shall be sweet. 3:25 Be not afraid of sudden fear, neither of the desolation sces et suauis erit somnus tuus 3:25 ne paueas repentino of the wicked, when it cometh.

3:26 For the LORD shall be thy confidence, and shall keep terrore et inruentes tibi potentias impiorum 3:26 Dominus

thy foot from being taken.

enim erit in latere tuo et custodiet pedem tuum ne capiaris

3:27 Withhold not good from them to whom it is due, 3:27 noli prohibere benefacere eum qui potest si uales et ipse

when it is in the power of thine hand to do it.

3:28 Say not unto thy neighbour, Go, and come again, and benefac 3:28 ne dicas amico tuo uade et reuertere et cras

to morrow I will give; when thou hast it by thee.

3:29 Devise not evil against thy neighbour, seeing he dabo tibi cum statim possis dare 3:29 ne moliaris amico tuo

dwelleth securely by thee.

3:30 Strive not with a man without cause, if he have done malum cum ille in te habeat fiduciam 3:30 ne contendas ad-

thee no harm.

3:31 Envy thou not the oppressor, and choose none of his uersus hominem frustra cum ipse tibi nihil mali fecerit 3:31

ways.

3:32 For the froward is abomination to the LORD: but his ne aemuleris hominem iniustum nec imiteris uias eius 3:32

secret is with the righteous.

quia abominatio Domini est omnis inlusor et cum simplici-

3:33 The curse of the LORD is in the house of the wicked: bus sermocinatio eius 3:33 egestas a Domino in domo impii

but he blesseth the habitation of the just.

3:34 Surely he scorneth the scorners: but he giveth grace habitacula autem iustorum benedicentur 3:34 inlusores ipse

unto the lowly.

3:35 The wise shall inherit glory: but shame shall be the deludet et mansuetis dabit gratiam 3:35 gloriam sapientes

promotion of fools.

possidebunt stultorum exaltatio ignominia

4:1 Hear, ye children, the instruction of a father, and attend 4:1 audite filii disciplinam patris et adtendite ut sciatis pru-

to know understanding. 4:2 For I give you good doctrine, forsake ye not my law. dentiam 4:2 donum bonum tribuam uobis legem meam ne

4:3 For I was my father’s son, tender and only beloved in derelinquatis 4:3 nam et ego filius fui patris mei tenellus

the sight of my mother.

4:4 He taught me also, and said unto me, Let thine heart et unigenitus coram matre mea 4:4 et docebat me atque di-

retain my words: keep my commandments, and live.

cebat suscipiat uerba mea cor tuum custodi praecepta mea

4:5 Get wisdom, get understanding: forget it not; neither et uiues 4:5 posside sapientiam posside prudentiam ne obdecline from the words of my mouth.

4:6 Forsake her not, and she shall preserve thee: love her, liuiscaris neque declines a uerbis oris mei 4:6 ne dimittas

and she shall keep thee.

4:7 Wisdom is the principal thing; therefore get wisdom: eam et custodiet te dilige eam et seruabit te 4:7 principium

and with all thy getting get understanding.

sapientiae posside sapientiam et in omni possessione tua ad-

4:8 Exalt her, and she shall promote thee: she shall bring quire prudentiam 4:8 arripe illam et exaltabit te glorifica-

thee to honour, when thou dost embrace her.

4:9 She shall give to thine head an ornament of grace: a beris ab ea cum eam fueris amplexatus 4:9 dabit capiti tuo

crown of glory shall she deliver to thee.

4:10 Hear, O my son, and receive my sayings; and the augmenta gratiarum et corona inclita proteget te 4:10 audi

years of thy life shall be many.

fili mi et suscipe uerba mea ut multiplicentur tibi anni ui-

4:11 I have taught thee in the way of wisdom; I have led tae 4:11 uiam sapientiae monstraui tibi duxi te per semi-

thee in right paths.

4:12 When thou goest, thy steps shall not be straitened; tas aequitatis 4:12 quas cum ingressus fueris non artabun-

and when thou runnest, thou shalt not stumble. 4:13 Take fast hold of instruction; let her not go: keep her; tur gressus tui et currens non habebis offendiculum 4:13 for she is thy life.

tene disciplinam ne dimittas eam custodi illam quia ipsa est

4:14 Enter not into the path of the wicked, and go not in uita tua 4:14 ne delecteris semitis impiorum nec tibi placeat

the way of evil men.

malorum uia 4:15 fuge ab ea ne transeas per illam declina 4:15 Avoid it, pass not by it, turn from it, and pass away. et desere eam 4:16 non enim dormiunt nisi malefecerint et 4:16 For they sleep not, except they have done mischief;

and their sleep is taken away, unless they cause some to fall.

rapitur somnus ab eis nisi subplantauerint 4:17 comedunt

4:17 For they eat the bread of wickedness, and drink the wine of violence.

panem impietatis et uinum iniquitatis bibunt 4:18 iustorum 4:18 But the path of the just is as the shining light, that

shineth more and more unto the perfect day.

autem semita quasi lux splendens procedit et crescit usque ad perfectam diem 4:19 uia impiorum tenebrosa nesciunt 4:19 The way of the wicked is as darkness: they know not

at what they stumble.

ubi corruant 4:20 fili mi ausculta sermones meos et ad elo- 4:20 My son, attend to my words; incline thine ear unto

my sayings.

quia mea inclina aurem tuam 4:21 ne recedant ab oculis tuis 4:21 Let them not depart from thine eyes; keep them in the

midst of thine heart.

custodi ea in medio cordis tui 4:22 uita enim sunt inuenien- 4:22 For they are life unto those that find them, and health

to all their flesh.

tibus ea et uniuersae carni sanitas 4:23 omni custodia serua 4:23 Keep thy heart with all diligence; for out of it are the

issues of life.

cor tuum quia ex ipso uita procedit 4:24 remoue a te os pra- 4:24 Put away from thee a froward mouth, and perverse

lips put far from thee.

uum et detrahentia labia sint procul a te 4:25 oculi tui re- 4:25 Let thine eyes look right on, and let thine eyelids look

straight before thee.

cta uideant et palpebrae tuae praecedant gressus tuos 4:26 4:26 Ponder the path of thy feet, and let all thy ways be

established.

dirige semitam pedibus tuis et omnes uiae tuae stabilientur 4:27 ne declines ad dexteram et ad sinistram auerte pedem 4:27 Turn not to the right hand nor to the left: remove thy

foot from evil.

tuum a malo

5:1 fili mi adtende sapientiam meam et prudentiae meae 5:1 My son, attend unto my wisdom, and bow thine ear to

my understanding:

inclina aurem tuam 5:2 ut custodias cogitationes et disci- 5:2 That thou mayest regard discretion, and that thy lips

may keep knowledge.

plinam labia tua conseruent 5:3 fauus enim stillans labia 5:3 For the lips of a strange woman drop as an honeycomb,

and her mouth is smoother than oil:

meretricis et nitidius oleo guttur eius 5:4 nouissima autem 5:4 But her end is bitter as wormwood, sharp as a two-

edged sword.

illius amara quasi absinthium et acuta quasi gladius biceps 5:5 pedes eius descendunt in mortem et ad inferos gressus il- 5:5 Her feet go down to death; her steps take hold on hell. lius penetrant 5:6 per semitam uitae non ambulat uagi sunt 5:6 Lest thou shouldest ponder the path of life, her ways

are moveable, that thou canst not know them.

gressus eius et inuestigabiles 5:7 nunc ergo fili audi me et 5:7 Hear me now therefore, O ye children, and depart not

from the words of my mouth.

ne recedas a uerbis oris mei 5:8 longe fac ab ea uiam tuam 5:8 Remove thy way far from her, and come not nigh the

door of her house:

et ne adpropinques foribus domus eius 5:9 ne des alienis 5:9 Lest thou give thine honour unto others, and thy years

unto the cruel:

honorem tuum et annos tuos crudeli 5:10 ne forte imple- 5:10 Lest strangers be filled with thy wealth; and thy

labours be in the house of a stranger;

antur extranei uiribus tuis et labores tui sint in domo aliena 5:11 et gemas in nouissimis quando consumpseris carnes et 5:11 And thou mourn at the last, when thy flesh and thy

body are consumed,

corpus tuum et dicas 5:12 cur detestatus sum disciplinam 5:12 And say, How have I hated instruction, and my heart

despised reproof;

et increpationibus non adquieuit cor meum 5:13 nec au- 5:13 And have not obeyed the voice of my teachers, nor

inclined mine ear to them that instructed me!

diui uocem docentium me et magistris non inclinaui aurem

meam 5:14 paene fui in omni malo in medio ecclesiae et 5:14 I was almost in all evil in the midst of the congregation and assembly.

5:15 Drink waters out of thine own cistern, and running synagogae 5:15 bibe aquam de cisterna tua et fluenta putei

waters out of thine own well.

5:16 Let thy fountains be dispersed abroad, and rivers of tui 5:16 deriuentur fontes tui foras et in plateis aquas tuas

waters in the streets.

5:17 Let them be only thine own, and not strangers’ with diuide 5:17 habeto eas solus nec sint alieni participes tui

thee.

5:18 Let thy fountain be blessed: and rejoice with the wife 5:18 sit uena tua benedicta et laetare cum muliere adulescen-

of thy youth.

5:19 Let her be as the loving hind and pleasant roe; let tiae tuae 5:19 cerua carissima et gratissimus hinulus ubera

her breasts satisfy thee at all times; and be thou ravished always with her love.

eius inebrient te omni tempore in amore illius delectare iu-

5:20 And why wilt thou, my son, be ravished with a giter 5:20 quare seduceris fili mi ab aliena et foueris sinu

strange woman, and embrace the bosom of a stranger?

5:21 For the ways of man are before the eyes of the LORD, alterius 5:21 respicit Dominus uias hominis et omnes gres-

and he pondereth all his goings.

5:22 His own iniquities shall take the wicked himself, and sus illius considerat 5:22 iniquitates suae capiunt impium et

he shall be holden with the cords of his sins. 5:23 He shall die without instruction; and in the greatness funibus peccatorum suorum constringitur 5:23 ipse morieof his folly he shall go astray.

tur quia non habuit disciplinam et multitudine stultitiae suae decipietur

6:1 My son, if thou be surety for thy friend, if thou hast 6:1 fili mi si spoponderis pro amico tuo defixisti apud ex-

stricken thy hand with a stranger,

6:2 Thou art snared with the words of thy mouth, thou art traneum manum tuam 6:2 inlaqueatus es uerbis oris tui et

taken with the words of thy mouth.

6:3 Do this now, my son, and deliver thyself, when thou captus propriis sermonibus 6:3 fac ergo quod dico fili mi

art come into the hand of thy friend; go, humble thyself, and make sure thy friend.

et temet ipsum libera quia incidisti in manu proximi tui di-

6:4 Give not sleep to thine eyes, nor slumber to thine eye- scurre festina suscita amicum tuum 6:4 ne dederis somlids.

6:5 Deliver thyself as a roe from the hand of the hunter, num oculis tuis nec dormitent palpebrae tuae 6:5 eruere

and as a bird from the hand of the fowler.

quasi dammula de manu et quasi auis de insidiis aucupis

6:6 Go to the ant, thou sluggard; consider her ways, and 6:6 uade ad formicam o piger et considera uias eius et disce

be wise:

6:7 Which having no guide, overseer, or ruler, sapientiam 6:7 quae cum non habeat ducem nec praecep-6:8 Provideth her meat in the summer, and gathereth her torem nec principem 6:8 parat aestate cibum sibi et con-

food in the harvest. 6:9 How long wilt thou sleep, O sluggard? when wilt thou gregat in messe quod comedat 6:9 usquequo piger dorarise out of thy sleep? 6:10 Yet a little sleep, a little slumber, a little folding of mis quando consurges ex somno tuo 6:10 paululum dorthe hands to sleep:

mies paululum dormitabis paululum conseres manus ut dor-

6:11 So shall thy poverty come as one that travelleth, and mias 6:11 et ueniet tibi quasi uiator egestas et pauperies

thy want as an armed man.

6:12 A naughty person, a wicked man, walketh with a quasi uir armatus 6:12 homo apostata uir inutilis graditur

froward mouth.

6:13 He winketh with his eyes, he speaketh with his feet, ore peruerso 6:13 annuit oculis terit pede digito loquitur

he teacheth with his fingers; 6:14 Frowardness is in his heart, he deviseth mischief con- 6:14 prauo corde machinatur malum et in omni tempore iurtinually; he soweth discord.

6:15 Therefore shall his calamity come suddenly; sud- gia seminat 6:15 huic extemplo ueniet perditio sua et su-

denly shall he be broken without remedy.

6:16 These six things doth the LORD hate: yea, seven are bito conteretur nec habebit ultra medicinam 6:16 sex sunt

an abomination unto him:

6:17 A proud look, a lying tongue, and hands that shed quae odit Dominus et septimum detestatur anima eius 6:17

innocent blood,

oculos sublimes linguam mendacem manus effundentes in-

noxium sanguinem 6:18 cor machinans cogitationes pes- 6:18 An heart that deviseth wicked imaginations, feet that

be swift in running to mischief,

simas pedes ueloces ad currendum in malum 6:19 profe- 6:19 A false witness that speaketh lies, and he that soweth

discord among brethren.

rentem mendacia testem fallacem et eum qui seminat inter fratres discordias 6:20 conserua fili mi praecepta patris tui 6:20 My son, keep thy father’s commandment, and forsake

not the law of thy mother:

et ne dimittas legem matris tuae 6:21 liga ea in corde tuo 6:21 Bind them continually upon thine heart, and tie them

about thy neck.

iugiter et circumda gutturi tuo 6:22 cum ambulaueris gra- 6:22 When thou goest, it shall lead thee; when thou sleep-

est, it shall keep thee; and when thou awakest, it shall talk with thee.

diantur tecum cum dormieris custodiant te et euigilans loquere cum eis 6:23 quia mandatum lucerna est et lex lux et 6:23 For the commandment is a lamp; and the law is light;

and reproofs of instruction are the way of life:

uia uitae increpatio disciplinae 6:24 ut custodiant te a mu- 6:24 To keep thee from the evil woman, from the flattery

of the tongue of a strange woman.

liere mala et a blanda lingua extraneae 6:25 non concupis- 6:25 Lust not after her beauty in thine heart; neither let her

take thee with her eyelids.

cat pulchritudinem eius cor tuum nec capiaris nutibus illius 6:26 pretium enim scorti uix unius est panis mulier autem 6:26 For by means of a whorish woman a man is brought

to a piece of bread: and the adultress will hunt for the precious life.

uiri pretiosam animam capit 6:27 numquid abscondere po-

6:27 Can a man take fire in his bosom, and his clothes not be burned?

test homo ignem in sinu suo ut uestimenta illius non ardeant 6:28 aut ambulare super prunas et non conburentur plantae 6:28 Can one go upon hot coals, and his feet not be

burned?

eius 6:29 sic qui ingreditur ad mulierem proximi sui non 6:29 So he that goeth in to his neighbour’s wife; whoso-

ever toucheth her shall not be innocent.

erit mundus cum tetigerit eam 6:30 non grandis est culpae 6:30 Men do not despise a thief, if he steal to satisfy his

soul when he is hungry;

cum quis furatus fuerit furatur enim ut esurientem impleat

animam 6:31 deprehensus quoque reddet septuplum et om- 6:31 But if he be found, he shall restore sevenfold; he shall

give all the substance of his house.

nem substantiam domus suae tradet 6:32 qui autem adulter 6:32 But whoso committeth adultery with a woman lack-

eth understanding: he that doeth it destroyeth his own soul.

est propter cordis inopiam perdet animam suam 6:33 turpi- 6:33 A wound and dishonour shall he get; and his reproach

shall not be wiped away.

tudinem et ignominiam congregat sibi et obprobrium illius non delebitur 6:34 quia zelus et furor uiri non parcet in die 6:34 For jealousy is the rage of a man: therefore he will

not spare in the day of vengeance.

uindictae 6:35 nec adquiescet cuiusquam precibus nec sus- 6:35 He will not regard any ransom; neither will he rest

content, though thou givest many gifts.

cipiet pro redeeptione dona plurima

7:1 fili mi custodi sermones meos et praecepta mea re- 7:1 My son, keep my words, and lay up my command-

ments with thee.

conde tibi 7:2 serua mandata mea et uiues et legem meam 7:2 Keep my commandments, and live; and my law as the

apple of thine eye.

quasi pupillam oculi tui 7:3 liga eam in digitis tuis scribe 7:3 Bind them upon thy fingers, write them upon the table

of thine heart.

illam in tabulis cordis tui 7:4 dic sapientiae soror mea es 7:4 Say unto wisdom, Thou art my sister; and call under-

standing thy kinswoman:

et prudentiam uoca amicam tuam 7:5 ut custodiat te a mu- 7:5 That they may keep thee from the strange woman,

from the stranger which flattereth with her words.

liere extranea et ab aliena quae uerba sua dulcia facit 7:6 7:6 For at the window of my house I looked through my

casement,

de fenestra enim domus meae per cancellos prospexi 7:7 et 7:7 And beheld among the simple ones, I discerned among

the youths, a young man void of understanding,

uideo paruulos considero uecordem iuuenem 7:8 qui trans- 7:8 Passing through the street near her corner; and he went

the way to her house,

it in platea iuxta angulum et propter uiam domus illius gra-

7:9 In the twilight, in the evening, in the black and dark ditur 7:9 in obscuro aduesperascente die in noctis tenebris

night:

7:10 And, behold, there met him a woman with the attire et caligine 7:10 et ecce mulier occurrit illi ornatu meret-

of an harlot, and subtil of heart.

7:11 (She is loud and stubborn; her feet abide not in her ricio praeparata ad capiendas animas garrula et uaga 7:11

house:

quietis inpatiens nec ualens in domo consistere pedibus suis

7:12 Now is she without, now in the streets, and lieth in 7:12 nunc foris nunc in plateis nunc iuxta angulos insidians

wait at every corner.)

7:13 So she caught him, and kissed him, and with an im- 7:13 adprehensumque deosculatur iuuenem et procaci uultu

pudent face said unto him,

7:14 I have peace offerings with me; this day have I payed blanditur dicens 7:14 uictimas pro salute debui hodie red-

my vows. 7:15 Therefore came I forth to meet thee, diligently to seek didi uota mea 7:15 idcirco egressa sum in occursum tuum thy face, and I have found thee.

7:16 I have decked my bed with coverings of tapestry, with desiderans te uidere et repperi 7:16 intexui funibus lectum

carved works, with fine linen of Egypt.

7:17 I have perfumed my bed with myrrh, aloes, and cin- meum straui tapetibus pictis ex Aegypto 7:17 aspersi cu-

namon. 7:18 Come, let us take our fill of love until the morning: bile meum murra et aloe et cinnamomo 7:18 ueni inebrielet us solace ourselves with loves.

mur uberibus donec inlucescat dies et fruamur cupitis am-

7:19 For the goodman is not at home, he is gone a long plexibus 7:19 non est enim uir in domo sua abiit uia lon-

journey:

7:20 He hath taken a bag of money with him, and will gissima 7:20 sacculum pecuniae secum tulit in die plenae

come home at the day appointed.

7:21 With her much fair speech she caused him to yield, lunae reuersurus est domum suam 7:21 inretiuit eum mul-

with the flattering of her lips she forced him.

7:22 He goeth after her straightway, as an ox goeth to the tis sermonibus et blanditiis labiorum protraxit illum 7:22

slaughter, or as a fool to the correction of the stocks;

statim eam sequitur quasi bos ductus ad uictimam et quasi agnus lasciuiens et ignorans quod ad uincula stultus trahatur

7:23 Till a dart strike through his liver; as a bird hasteth to 7:23 donec transfigat sagitta iecur eius uelut si auis festinet ad

the snare, and knoweth not that it is for his life.

7:24 Hearken unto me now therefore, O ye children, and laqueum et nescit quia de periculo animae illius agitur 7:24

attend to the words of my mouth. 7:25 Let not thine heart decline to her ways, go not astray nunc ergo fili audi me et adtende uerba oris mei 7:25 ne abin her paths.

strahatur in uiis illius mens tua neque decipiaris semitis eius

7:26 For she hath cast down many wounded: yea, many 7:26 multos enim uulneratos deiecit et fortissimi quique in-

strong men have been slain by her.

7:27 Her house is the way to hell, going down to the cham- terfecti sunt ab ea 7:27 uiae inferi domus eius penetrantes

bers of death.

interiora mortis

8:1 Doth not wisdom cry? and understanding put forth her 8:1 numquid non sapientia clamitat et prudentia dat uovoice?

8:2 She standeth in the top of high places, by the way in cem suam 8:2 in summis excelsisque uerticibus super uiam

the places of the paths.

8:3 She crieth at the gates, at the entry of the city, at the in mediis semitis stans 8:3 iuxta portas ciuitatis in ipsis fo-

coming in at the doors.

8:4 Unto you, O men, I call; and my voice is to the sons of ribus loquitur dicens 8:4 o uiri ad uos clamito et uox mea

man.

8:5 O ye simple, understand wisdom: and, ye fools, be ye ad filios hominum 8:5 intellegite paruuli astutiam et insi-

of an understanding heart. 8:6 Hear; for I will speak of excellent things; and the open- pientes animaduertite 8:6 audite quoniam de rebus magnis ing of my lips shall be right things.

8:7 For my mouth shall speak truth; and wickedness is an locutura sum et aperientur labia mea ut recta praedicent 8:7

abomination to my lips.

ueritatem meditabitur guttur meum et labia mea detestabuntur impium 8:8 iusti sunt omnes sermones mei non est in eis 8:8 All the words of my mouth are in righteousness; there

is nothing froward or perverse in them.

prauum quid neque peruersum 8:9 recti sunt intellegentibus 8:9 They are all plain to him that understandeth, and right

to them that find knowledge.

et aequi inuenientibus scientiam 8:10 accipite disciplinam 8:10 Receive my instruction, and not silver; and knowl-

edge rather than choice gold.

meam et non pecuniam doctrinam magis quam aurum eligite 8:11 melior est enim sapientia cunctis pretiosissimis et omne 8:11 For wisdom is better than rubies; and all the things

that may be desired are not to be compared to it.

desiderabile ei non potest conparari 8:12 ego sapientia ha- 8:12 I wisdom dwell with prudence, and find out knowl-

edge of witty inventions.

bito in consilio et eruditis intersum cogitationibus 8:13 ti- 8:13 The fear of the LORD is to hate evil: pride, and ar-

rogancy, and the evil way, and the froward mouth, do I hate.

mor Domini odit malum arrogantiam et superbiam et uiam prauam et os bilingue detestor 8:14 meum est consilium et 8:14 Counsel is mine, and sound wisdom: I am under-

standing; I have strength.

aequitas mea prudentia mea est fortitudo 8:15 per me reges 8:15 By me kings reign, and princes decree justice.

regnant et legum conditores iusta decernunt 8:16 per me 8:16 By me princes rule, and nobles, even all the judges of

the earth.

principes imperant et potentes decernunt iustitiam 8:17 ego 8:17 I love them that love me; and those that seek me early

shall find me.

diligentes me diligo et qui mane uigilant ad me inuenient me 8:18 mecum sunt diuitiae et gloria opes superbae et iustitia 8:18 Riches and honour are with me; yea, durable riches

and righteousness.

8:19 melior est fructus meus auro et pretioso lapide et geni- 8:19 My fruit is better than gold, yea, than fine gold; and

my revenue than choice silver.

mina mea argento electo 8:20 in uiis iustitiae ambulo in me- 8:20 I lead in the way of righteousness, in the midst of the

paths of judgment:

dio semitarum iudicii 8:21 ut ditem diligentes me et thesau- 8:21 That I may cause those that love me to inherit sub-

stance; and I will fill their treasures.

ros eorum repleam 8:22 Dominus possedit me initium uia- 8:22 The LORD possessed me in the beginning of his way,

before his works of old.

rum suarum antequam quicquam faceret a principio 8:23 8:23 I was set up from everlasting, from the beginning, or

ever the earth was.

ab aeterno ordita sum et ex antiquis antequam terra fieret 8:24 necdum erant abyssi et ego iam concepta eram necdum 8:24 When there were no depths, I was brought forth;

when there were no fountains abounding with water.

fontes aquarum eruperant 8:25 necdum montes graui mole 8:25 Before the mountains were settled, before the hills

was I brought forth:

constiterant ante colles ego parturiebar 8:26 adhuc terram 8:26 While as yet he had not made the earth, nor the fields,

nor the highest part of the dust of the world.

non fecerat et flumina et cardines orbis terrae 8:27 quando 8:27 When he prepared the heavens, I was there: when he

set a compass upon the face of the depth:

praeparabat caelos aderam quando certa lege et gyro uallabat abyssos 8:28 quando aethera firmabat sursum et libra- 8:28 When he established the clouds above: when he

strengthened the fountains of the deep:

bat fontes aquarum 8:29 quando circumdabat mari termi- 8:29 When he gave to the sea his decree, that the waters

should not pass his commandment: when he appointed the foundations of the earth:

num suum et legem ponebat aquis ne transirent fines suos quando adpendebat fundamenta terrae 8:30 cum eo eram 8:30 Then I was by him, as one brought up with him: and

I was daily his delight, rejoicing always before him;

cuncta conponens et delectabar per singulos dies ludens co-

ram eo omni tempore 8:31 ludens in orbe terrarum et de- 8:31 Rejoicing in the habitable part of his earth; and my

delights were with the sons of men.

liciae meae esse cum filiis hominum 8:32 nunc ergo filii 8:32 Now therefore hearken unto me, O ye children: for

blessed are they that keep my ways.

8:33 Hear instruction, and be wise, and refuse it not. audite me beati qui custodiunt uias meas 8:33 audite disci-8:34 Blessed is the man that heareth me, watching daily at plinam et estote sapientes et nolite abicere eam 8:34 beatus

my gates, waiting at the posts of my doors.

homo qui audit me qui uigilat ad fores meas cotidie et obser-

8:35 For whoso findeth me findeth life, and shall obtain uat ad postes ostii mei 8:35 qui me inuenerit inueniet uitam

favour of the LORD. 8:36 But he that sinneth against me wrongeth his own soul: et hauriet salutem a Domino 8:36 qui autem in me peccaall they that hate me love death.

uerit laedet animam suam omnes qui me oderunt diligunt mortem

9:1 Wisdom hath builded her house, she hath hewn out her 9:1 sapientia aedificauit sibi domum excidit columnas se-

seven pillars: 9:2 She hath killed her beasts; she hath mingled her wine; ptem 9:2 immolauit uictimas suas miscuit uinum et proposhe hath also furnished her table. 9:3 She hath sent forth her maidens: she crieth upon the suit mensam suam 9:3 misit ancillas suas ut uocarent ad highest places of the city,

9:4 Whoso is simple, let him turn in hither: as for him that arcem et ad moenia ciuitatis 9:4 si quis est paruulus ueniat

wanteth understanding, she saith to him,

9:5 Come, eat of my bread, and drink of the wine which I ad me et insipientibus locuta est 9:5 uenite comedite panem

have mingled.

9:6 Forsake the foolish, and live; and go in the way of meum et bibite uinum quod miscui uobis 9:6 relinquite in-

understanding. 9:7 He that reproveth a scorner getteth to himself shame: fantiam et uiuite et ambulate per uias prudentiae 9:7 qui and he that rebuketh a wicked man getteth himself a blot.

erudit derisorem ipse sibi facit iniuriam et qui arguit impium

9:8 Reprove not a scorner, lest he hate thee: rebuke a wise generat maculam sibi 9:8 noli arguere derisorem ne oderit

man, and he will love thee.

9:9 Give instruction to a wise man, and he will be yet te argue sapientem et diliget te 9:9 da sapienti et addetur

wiser: teach a just man, and he will increase in learning.

9:10 The fear of the LORD is the beginning of wisdom: ei sapientia doce iustum et festinabit accipere 9:10 princi-

and the knowledge of the holy is understanding.

pium sapientiae timor Domini et scientia sanctorum pruden-

9:11 For by me thy days shall be multiplied, and the years tia 9:11 per me enim multiplicabuntur dies tui et addentur

of thy life shall be increased.

9:12 If thou be wise, thou shalt be wise for thyself: but if tibi anni uitae 9:12 si sapiens fueris tibimet ipsi eris si in-

thou scornest, thou alone shalt bear it.

9:13 A foolish woman is clamorous: she is simple, and lusor solus portabis malum 9:13 mulier stulta et clamosa

knoweth nothing.

9:14 For she sitteth at the door of her house, on a seat in plenaque inlecebris et nihil omnino sciens 9:14 sedit in fo-

the high places of the city,

9:15 To call passengers who go right on their ways: ribus domus suae super sellam in excelso urbis loco 9:15 ut 9:16 Whoso is simple, let him turn in hither: and as for uocaret transeuntes uiam et pergentes itinere suo 9:16 quis

him that wanteth understanding, she saith to him,

9:17 Stolen waters are sweet, and bread eaten in secret is est paruulus declinet ad me et uecordi locuta est 9:17 aquae

pleasant.

9:18 But he knoweth not that the dead are there; and that furtiuae dulciores sunt et panis absconditus suauior 9:18 et

her guests are in the depths of hell.

ignorauit quod gigantes ibi sint et in profundis inferni conuiuae eius

10:1 The proverbs of Solomon. A wise son maketh a glad 10:1 parabolae Salomonis filius sapiens laetificat patrem

father: but a foolish son is the heaviness of his mother.

10:2 Treasures of wickedness profit nothing: but righ- filius uero stultus maestitia est matris suae 10:2 non prode-

teousness delivereth from death.

10:3 The LORD will not suffer the soul of the righteous to eunt thesauri impietatis iustitia uero liberabit a morte 10:3

famish: but he casteth away the substance of the wicked.

non adfliget Dominus fame animam iusti et insidias impiorum subuertet 10:4 egestatem operata est manus remissa 10:4 He becometh poor that dealeth with a slack hand: but

the hand of the diligent maketh rich.

manus autem fortium diuitias parat 10:5 qui congregat in 10:5 He that gathereth in summer is a wise son: but he that

sleepeth in harvest is a son that causeth shame.

messe filius sapiens est qui autem stertit aestate filius confusionis 10:6 benedictio super caput iusti os autem impiorum 10:6 Blessings are upon the head of the just: but violence

covereth the mouth of the wicked.

operit iniquitatem 10:7 memoria iusti cum laudibus et no- 10:7 The memory of the just is blessed: but the name of

the wicked shall rot.

men impiorum putrescet 10:8 sapiens corde praecepta su- 10:8 The wise in heart will receive commandments: but a

prating fool shall fall.

scipiet stultus caeditur labiis 10:9 qui ambulat simpliciter 10:9 He that walketh uprightly walketh surely: but he that

perverteth his ways shall be known.

ambulat confidenter qui autem deprauat uias suas manifestus erit 10:10 qui annuit oculo dabit dolorem stultus labiis uer- 10:10 He that winketh with the eye causeth sorrow: but a

prating fool shall fall.

berabitur 10:11 uena uitae os iusti et os impiorum operiet 10:11 The mouth of a righteous man is a well of life: but

violence covereth the mouth of the wicked.

iniquitatem 10:12 odium suscitat rixas et uniuersa delicta 10:12 Hatred stirreth up strifes: but love covereth all sins.

operit caritas 10:13 in labiis sapientis inuenietur sapientia 10:13 In the lips of him that hath understanding wisdom

is found: but a rod is for the back of him that is void of understanding.

et uirga in dorso eius qui indiget corde 10:14 sapientes ab-

10:14 Wise men lay up knowledge: but the mouth of the foolish is near destruction.

scondunt scientiam os autem stulti confusioni proximum est 10:15 substantia diuitis urbs fortitudinis eius pauor paupe- 10:15 The rich man’s wealth is his strong city: the destruc-

tion of the poor is their poverty.

rum egestas eorum 10:16 opus iusti ad uitam fructus impii 10:16 The labour of the righteous tendeth to life: the fruit

of the wicked to sin.

ad peccatum 10:17 uia uitae custodienti disciplinam qui au- 10:17 He is in the way of life that keepeth instruction: but

he that refuseth reproof erreth.

tem increpationes relinquit errat 10:18 abscondunt odium 10:18 He that hideth hatred with lying lips, and he that

uttereth a slander, is a fool.

labia mendacia qui profert contumeliam insipiens est 10:19 10:19 In the multitude of words there wanteth not sin: but

he that refraineth his lips is wise.

in multiloquio peccatum non deerit qui autem moderatur labia sua prudentissimus est 10:20 argentum electum lingua 10:20 The tongue of the just is as choice silver: the heart

of the wicked is little worth.

iusti cor impiorum pro nihilo 10:21 labia iusti erudiunt plu- 10:21 The lips of the righteous feed many: but fools die

for want of wisdom.

rimos qui autem indocti sunt in cordis egestate morientur 10:22 benedictio Domini diuites facit nec sociabitur ei ad- 10:22 The blessing of the LORD, it maketh rich, and he

addeth no sorrow with it.

flictio 10:23 quasi per risum stultus operatur scelus sapien- 10:23 It is as sport to a fool to do mischief: but a man of

understanding hath wisdom.

tia autem est uiro prudentia 10:24 quod timet impius ueniet 10:24 The fear of the wicked, it shall come upon him: but

the desire of the righteous shall be granted.

super eum desiderium suum iustis dabitur 10:25 quasi tem- 10:25 As the whirlwind passeth, so is the wicked no more:

but the righteous is an everlasting foundation.

pestas transiens non erit impius iustus autem quasi fundamentum sempiternum 10:26 sicut acetum dentibus et fumus 10:26 As vinegar to the teeth, and as smoke to the eyes, so

is the sluggard to them that send him.

oculis sic piger his qui miserunt eum 10:27 timor Domini 10:27 The fear of the LORD prolongeth days: but the

years of the wicked shall be shortened.

adponet dies et anni impiorum breuiabuntur 10:28 expec- 10:28 The hope of the righteous shall be gladness: but the

expectation of the wicked shall perish.

tatio iustorum laetitia spes autem impiorum peribit 10:29 10:29 The way of the LORD is strength to the upright: but

destruction shall be to the workers of iniquity.

fortitudo simplicis uia Domini et pauor his qui operantur 10:30 The righteous shall never be removed: but the malum 10:30 iustus in aeternum non commouebitur impii

wicked shall not inhabit the earth.

10:31 The mouth of the just bringeth forth wisdom: but autem non habitabunt in terram 10:31 os iusti parturiet sa-

the froward tongue shall be cut out.

10:32 The lips of the righteous know what is acceptable: pientiam lingua prauorum peribit 10:32 labia iusti conside-

but the mouth of the wicked speaketh frowardness.

rant placita et os impiorum peruersa

11:1 A false balance is abomination to the LORD: but a 11:1 statera dolosa abominatio apud Dominum et pondus

just weight is his delight.

11:2 When pride cometh, then cometh shame: but with the aequum uoluntas eius 11:2 ubi fuerit superbia ibi erit et

lowly is wisdom. 11:3 The integrity of the upright shall guide them: but the contumelia ubi autem humilitas ibi et sapientia 11:3 simperverseness of transgressors shall destroy them.

plicitas iustorum diriget eos et subplantatio peruersorum ua-

11:4 Riches profit not in the day of wrath: but righteous- stabit illos 11:4 non prodeeunt diuitiae in die ultionis iu-

ness delivereth from death.

11:5 The righteousness of the perfect shall direct his way: stitia autem liberabit a morte 11:5 iustitia simplicis diri-

but the wicked shall fall by his own wickedness. 11:6 The righteousness of the upright shall deliver them: get uiam eius et in impietate sua corruet impius 11:6 iusbut transgressors shall be taken in their own naughtiness.

titia rectorum liberabit eos et in insidiis suis capientur iniqui

11:7 When a wicked man dieth, his expectation shall per- 11:7 mortuo homine impio nulla erit ultra spes et expectatio

ish: and the hope of unjust men perisheth.

11:8 The righteous is delivered out of trouble, and the sollicitorum peribit 11:8 iustus de angustia liberatus est et

wicked cometh in his stead.

11:9 An hypocrite with his mouth destroyeth his neigh- tradetur impius pro eo 11:9 simulator ore decipit amicum

bour: but through knowledge shall the just be delivered.

11:10 When it goeth well with the righteous, the city re- suum iusti autem liberabuntur scientia 11:10 in bonis iusto-

joiceth: and when the wicked perish, there is shouting.

rum exultabit ciuitas et in perditione impiorum erit laudatio

11:11 By the blessing of the upright the city is exalted: but 11:11 benedictione iustorum exaltabitur ciuitas et ore impio-

it is overthrown by the mouth of the wicked.

11:12 He that is void of wisdom despiseth his neighbour: rum subuertetur 11:12 qui despicit amicum suum indigens

but a man of understanding holdeth his peace.

11:13 A talebearer revealeth secrets: but he that is of a corde est uir autem prudens tacebit 11:13 qui ambulat frau-

faithful spirit concealeth the matter.

dulenter reuelat arcana qui autem fidelis est animi celat com-

11:14 Where no counsel is, the people fall: but in the mul- missum 11:14 ubi non est gubernator populus corruet satitude of counsellors there is safety. 11:15 He that is surety for a stranger shall smart for it: and lus autem ubi multa consilia 11:15 adfligetur malo qui fihe that hateth suretiship is sure.

dem facit pro extraneo qui autem cauet laqueos securus erit

11:16 A gracious woman retaineth honour: and strong 11:16 mulier gratiosa inueniet gloriam et robusti habebunt di-

men retain riches.

11:17 The merciful man doeth good to his own soul: but uitias 11:17 benefacit animae suae uir misericors qui autem

he that is cruel troubleth his own flesh.

11:18 The wicked worketh a deceitful work: but to him crudelis est et propinquos abicit 11:18 impius facit opus in-

that soweth righteousness shall be a sure reward.

11:19 As righteousness tendeth to life: so he that pursueth stabile seminanti autem iustitiam merces fidelis 11:19 cle-

evil pursueth it to his own death.

11:20 They that are of a froward heart are abomination to mentia praeparat uitam et sectatio malorum mortem 11:20

the LORD: but such as are upright in their way are his delight.

abominabile Domino prauum cor et uoluntas eius in his qui

11:21 Though hand join in hand, the wicked shall not be simpliciter ambulant 11:21 manus in manu non erit inno-

unpunished: but the seed of the righteous shall be delivered.

cens malus semen autem iustorum saluabitur 11:22 circulus

11:22 As a jewel of gold in a swine’s snout, so is a fair woman which is without discretion.

11:23 The desire of the righteous is only good: but the aureus in naribus suis mulier pulchra et fatua 11:23 deside-

expectation of the wicked is wrath.

rium iustorum omne bonum est praestolatio impiorum furor 11:24 alii diuidunt propria et ditiores fiunt alii rapiunt non 11:24 There is that scattereth, and yet increaseth; and there

is that withholdeth more than is meet, but it tendeth to poverty.

sua et semper in egestate sunt 11:25 anima quae benedicit

11:25 The liberal soul shall be made fat: and he that watereth shall be watered also himself.

inpinguabitur et qui inebriat ipse quoque inebriabitur 11:26 11:26 He that withholdeth corn, the people shall curse

him: but blessing shall be upon the head of him that selleth it.

qui abscondit frumenta maledicetur in populis benedictio autem super caput uendentium 11:27 bene consurgit diluculo 11:27 He that diligently seeketh good procureth favour:

but he that seeketh mischief, it shall come unto him.

qui quaerit bona qui autem inuestigator malorum est opprimetur ab eis 11:28 qui confidet in diuitiis suis corruet iusti 11:28 He that trusteth in his riches shall fall; but the righ-

teous shall flourish as a branch.

autem quasi uirens folium germinabunt 11:29 qui contur- 11:29 He that troubleth his own house shall inherit the

wind: and the fool shall be servant to the wise of heart.

bat domum suam possidebit uentos et qui stultus est seruiet sapienti 11:30 fructus iusti lignum uitae et qui suscipit ani- 11:30 The fruit of the righteous is a tree of life; and he that

winneth souls is wise.

mas sapiens est 11:31 si iustus in terra recipit quanto magis 11:31 Behold, the righteous shall be recompensed in the

earth: much more the wicked and the sinner.

impius et peccator

12:1 qui diligit disciplinam diligit scientiam qui autem odit 12:1 Whoso loveth instruction loveth knowledge: but he

that hateth reproof is brutish.

increpationes insipiens est 12:2 qui bonus est hauriet a Do- 12:2 A good man obtaineth favour of the LORD: but a

man of wicked devices will he condemn.

mino gratiam qui autem confidit cogitationibus suis impie agit 12:3 non roborabitur homo ex impietate et radix ius- 12:3 A man shall not be established by wickedness: but

the root of the righteous shall not be moved.

torum non commouebitur 12:4 mulier diligens corona uiro 12:4 A virtuous woman is a crown to her husband: but she

that maketh ashamed is as rottenness in his bones.

suo et putredo in ossibus eius quae confusione res dignas gerit 12:5 cogitationes iustorum iudicia et consilia impio- 12:5 The thoughts of the righteous are right: but the coun-

sels of the wicked are deceit.

rum fraudulentia 12:6 uerba impiorum insidiantur sanguini 12:6 The words of the wicked are to lie in wait for blood:

but the mouth of the upright shall deliver them.

os iustorum liberabit eos 12:7 uerte impios et non erunt do- 12:7 The wicked are overthrown, and are not: but the

house of the righteous shall stand.

mus autem iustorum permanebit 12:8 doctrina sua noscetur 12:8 A man shall be commended according to his wisdom:

but he that is of a perverse heart shall be despised.

uir qui autem uanus et excors est patebit contemptui 12:9 12:9 He that is despised, and hath a servant, is better than

he that honoureth himself, and lacketh bread.

melior est pauper et sufficiens sibi quam gloriosus et indi-

gens pane 12:10 nouit iustus animas iumentorum suorum 12:10 A righteous man regardeth the life of his beast: but

the tender mercies of the wicked are cruel.

uiscera autem impiorum crudelia 12:11 qui operatur terram 12:11 He that tilleth his land shall be satisfied with bread:

but he that followeth vain persons is void of understanding.

suam saturabitur panibus qui autem sectatur otium stultissi-

mus est 12:12 desiderium impii munimentum est pessimo- 12:12 The wicked desireth the net of evil men: but the root

of the righteous yieldeth fruit.

rum radix autem iustorum proficiet 12:13 propter peccata 12:13 The wicked is snared by the transgression of his lips:

but the just shall come out of trouble.

labiorum ruina proximat malo effugiet autem iustus de angustia 12:14 de fructu oris sui unusquisque replebitur bo- 12:14 A man shall be satisfied with good by the fruit of

his mouth: and the recompence of a man’s hands shall be rendered unto him.

nis et iuxta opera manuum suarum retribuetur ei 12:15 uia

12:15 The way of a fool is right in his own eyes: but he that hearkeneth unto counsel is wise.

stulti recta in oculis eius qui autem sapiens est audit con-

12:16 A fool’s wrath is presently known: but a prudent silia 12:16 fatuus statim indicat iram suam qui autem dis-

man covereth shame.

12:17 He that speaketh truth sheweth forth righteousness: simulat iniuriam callidus est 12:17 qui quod nouit loquitur

but a false witness deceit.

index iustitiae est qui autem mentitur testis est fraudulentus

12:18 There is that speaketh like the piercings of a sword: 12:18 est qui promittit et quasi gladio pungitur conscientiae

but the tongue of the wise is health.

12:19 The lip of truth shall be established for ever: but a lingua autem sapientium sanitas est 12:19 labium ueritatis

lying tongue is but for a moment.

firmum erit in perpetuum qui autem testis est repentinus con-

12:20 Deceit is in the heart of them that imagine evil: but cinnat linguam mendacii 12:20 dolus in corde cogitantium

to the counsellors of peace is joy.

mala qui autem ineunt pacis consilia sequitur eos gaudium

12:21 There shall no evil happen to the just: but the wicked 12:21 non contristabit iustum quicquid ei acciderit impii au-

shall be filled with mischief.

12:22 Lying lips are abomination to the LORD: but they tem replebuntur malo 12:22 abominatio Domino labia men-

that deal truly are his delight.

12:23 A prudent man concealeth knowledge: but the heart dacia qui autem fideliter agunt placent ei 12:23 homo uer-

of fools proclaimeth foolishness.

sutus celat scientiam et cor insipientium prouocabit stulti-

12:24 The hand of the diligent shall bear rule: but the tiam 12:24 manus fortium dominabitur quae autem remissa

slothful shall be under tribute.

12:25 Heaviness in the heart of man maketh it stoop: but est tributis seruiet 12:25 maeror in corde uiri humiliabit il-

a good word maketh it glad. 12:26 The righteous is more excellent than his neighbour: lud et sermone bono laetificabitur 12:26 qui neglegit dambut the way of the wicked seduceth them.

num propter amicum iustus est iter autem impiorum deci-

12:27 The slothful man roasteth not that which he took in piet eos 12:27 non inueniet fraudulentus lucrum et substan-

hunting: but the substance of a diligent man is precious.

12:28 In the way of righteousness is life: and in the path- tia hominis erit auri pretium 12:28 in semita iustitiae uita

way thereof there is no death.

iter autem deuium ducit ad mortem

13:1 A wise son heareth his father’s instruction: but a 13:1 filius sapiens doctrina patris qui autem inlusor est non

scorner heareth not rebuke.

13:2 A man shall eat good by the fruit of his mouth: but audit cum arguitur 13:2 de fructu oris homo saturabitur bo-

the soul of the transgressors shall eat violence.

13:3 He that keepeth his mouth keepeth his life: but he nis anima autem praeuaricatorum iniqua 13:3 qui custodit

that openeth wide his lips shall have destruction.

os suum custodit animam suam qui autem inconsideratus est

13:4 The soul of the sluggard desireth, and hath nothing: ad loquendum sentiet mala 13:4 uult et non uult piger anbut the soul of the diligent shall be made fat.

13:5 A righteous man hateth lying: but a wicked man is ima autem operantium inpinguabitur 13:5 uerbum mendax

loathsome, and cometh to shame.

13:6 Righteousness keepeth him that is upright in the way: iustus detestabitur impius confundit et confundetur 13:6 iu-

but wickedness overthroweth the sinner.

stitia custodit innocentis uiam impietas uero peccato sub-

13:7 There is that maketh himself rich, yet hath nothing: plantat 13:7 est quasi diues cum nihil habeat et est quasi there is that maketh himself poor, yet hath great riches.

13:8 The ransom of a man’s life are his riches: but the poor pauper cum in multis diuitiis sit 13:8 redeeptio animae uiri

heareth not rebuke.

diuitiae suae qui autem pauper est increpationem non sus-

13:9 The light of the righteous rejoiceth: but the lamp of tinet 13:9 lux iustorum laetificat lucerna autem impiorum

the wicked shall be put out.

13:10 Only by pride cometh contention: but with the well extinguetur 13:10 inter superbos semper iurgia sunt qui au-

advised is wisdom.

tem agunt cuncta consilio reguntur sapientia 13:11 substan- 13:11 Wealth gotten by vanity shall be diminished: but he

that gathereth by labour shall increase.

tia festinata minuetur quae autem paulatim colligitur manu multiplicabitur 13:12 spes quae differtur adfligit animam 13:12 Hope deferred maketh the heart sick: but when the

desire cometh, it is a tree of life.

lignum uitae desiderium ueniens 13:13 qui detrahit alicui 13:13 Whoso despiseth the word shall be destroyed: but

he that feareth the commandment shall be rewarded.

rei ipse se in futurum obligat qui autem timet praeceptum in pace uersabitur 13:14 lex sapientis fons uitae ut decli- 13:14 The law of the wise is a fountain of life, to depart

from the snares of death.

net a ruina mortis 13:15 doctrina bona dabit gratiam in iti- 13:15 Good understanding giveth favour: but the way of

transgressors is hard.

nere contemptorum uorago 13:16 astutus omnia agit cum 13:16 Every prudent man dealeth with knowledge: but a

fool layeth open his folly.

consilio qui autem fatuus est aperit stultitiam 13:17 nuntius 13:17 A wicked messenger falleth into mischief: but a

faithful ambassador is health.

impii cadet in malum legatus fidelis sanitas 13:18 egestas et 13:18 Poverty and shame shall be to him that refuseth in-

struction: but he that regardeth reproof shall be honoured.

ignominia ei qui deserit disciplinam qui autem adquiescit arguenti glorificabitur 13:19 desiderium si conpleatur delec- 13:19 The desire accomplished is sweet to the soul: but it

is abomination to fools to depart from evil.

tat animam detestantur stulti eos qui fugiunt mala 13:20 qui 13:20 He that walketh with wise men shall be wise: but a

companion of fools shall be destroyed.

cum sapientibus graditur sapiens erit amicus stultorum efficietur similis 13:21 peccatores persequetur malum et iustis 13:21 Evil pursueth sinners: but to the righteous good

shall be repayed.

retribuentur bona 13:22 bonus relinquet heredes filios et ne- 13:22 A good man leaveth an inheritance to his children’s

children: and the wealth of the sinner is laid up for the just.

potes et custoditur iusto substantia peccatoris 13:23 multi

13:23 Much food is in the tillage of the poor: but there is that is destroyed for want of judgment.

cibi in noualibus patrum et alii congregantur absque iudicio 13:24 qui parcit uirgae suae odit filium suum qui autem diligit 13:24 He that spareth his rod hateth his son: but he that

loveth him chasteneth him betimes.

illum instanter erudit 13:25 iustus comedit et replet animam 13:25 The righteous eateth to the satisfying of his soul:

but the belly of the wicked shall want.

suam uenter autem impiorum insaturabilis

14:1 sapiens mulier aedificauit domum suam insipiens in- 14:1 Every wise woman buildeth her house: but the foolish

plucketh it down with her hands.

structam quoque destruet manibus 14:2 ambulans recto iti- 14:2 He that walketh in his uprightness feareth the LORD:

but he that is perverse in his ways despiseth him.

nere et timens Deum despicitur ab eo qui infami graditur uia 14:3 in ore stulti uirga superbiae labia sapientium custo- 14:3 In the mouth of the foolish is a rod of pride: but the

lips of the wise shall preserve them.

diunt eos 14:4 ubi non sunt boues praesepe uacuum est ubi 14:4 Where no oxen are, the crib is clean: but much in-

crease is by the strength of the ox.

autem plurimae segetes ibi manifesta fortitudo bouis 14:5 14:5 A faithful witness will not lie: but a false witness will

utter lies.

testis fidelis non mentietur profert mendacium testis dolosus 14:6 quaerit derisor sapientiam et non inueniet doct- 14:6 A scorner seeketh wisdom, and findeth it not: but

knowledge is easy unto him that understandeth.

rina prudentium facilis 14:7 uade contra uirum stultum et 14:7 Go from the presence of a foolish man, when thou

perceivest not in him the lips of knowledge.

nescito labia prudentiae 14:8 sapientia callidi est intelle- 14:8 The wisdom of the prudent is to understand his way:

but the folly of fools is deceit.

gere uiam suam et inprudentia stultorum errans 14:9 stul- 14:9 Fools make a mock at sin: but among the righteous

there is favour.

tis inludet peccatum inter iustos morabitur gratia 14:10 cor 14:10 The heart knoweth his own bitterness; and a stranger

doth not intermeddle with his joy.

quod nouit amaritudinem animae suae in gaudio eius non

14:11 The house of the wicked shall be overthrown: but miscebitur extraneus 14:11 domus impiorum delebitur ta-

the tabernacle of the upright shall flourish.

14:12 There is a way which seemeth right unto a man, but bernacula iustorum germinabunt 14:12 est uia quae uide-

the end thereof are the ways of death.

tur homini iusta nouissima autem eius deducunt ad mor-

14:13 Even in laughter the heart is sorrowful; and the end tem 14:13 risus dolore miscebitur et extrema gaudii luctus

of that mirth is heaviness.

14:14 The backslider in heart shall be filled with his own occupat 14:14 uiis suis replebitur stultus et super eum erit

ways: and a good man shall be satisfied from himself.

14:15 The simple believeth every word: but the prudent uir bonus 14:15 innocens credit omni uerbo astutus consi-

man looketh well to his going. 14:16 A wise man feareth, and departeth from evil: but the derat gressus suos 14:16 sapiens timet et declinat malum fool rageth, and is confident.

14:17 He that is soon angry dealeth foolishly: and a man stultus transilit et confidit 14:17 inpatiens operabitur stul-

of wicked devices is hated.

14:18 The simple inherit folly: but the prudent are titiam et uir uersutus odiosus est 14:18 possidebunt par-

crowned with knowledge.

14:19 The evil bow before the good; and the wicked at the uuli stultitiam et astuti expectabunt scientiam 14:19 iace-

gates of the righteous.

14:20 The poor is hated even of his own neighbour: but bunt mali ante bonos et impii ante portas iustorum 14:20

the rich hath many friends.

etiam proximo suo pauper odiosus erit amici uero diuitum

14:21 He that despiseth his neighbour sinneth: but he that multi 14:21 qui despicit proximum suum peccat qui autem hath mercy on the poor, happy is he.

14:22 Do they not err that devise evil? but mercy and truth miseretur pauperi beatus erit 14:22 errant qui operantur ma-

shall be to them that devise good.

14:23 In all labour there is profit: but the talk of the lips lum misericordia et ueritas praeparant bona 14:23 in omni

tendeth only to penury.

opere erit abundantia ubi autem uerba sunt plurima frequen-

14:24 The crown of the wise is their riches: but the fool- ter egestas 14:24 corona sapientium diuitiae eorum fatuitas

ishness of fools is folly.

14:25 A true witness delivereth souls: but a deceitful wit- stultorum inprudentia 14:25 liberat animas testis fidelis et

ness speaketh lies.

14:26 In the fear of the LORD is strong confidence: and profert mendacia uersipellis 14:26 in timore Domini fidu-

his children shall have a place of refuge.

14:27 The fear of the LORD is a fountain of life, to depart cia fortitudinis et filiis eius erit spes 14:27 timor Domini

from the snares of death.

14:28 In the multitude of people is the king’s honour: but fons uitae ut declinet a ruina mortis 14:28 in multitudine

in the want of people is the destruction of the prince.

populi dignitas regis et in paucitate plebis ignominia prin-

14:29 He that is slow to wrath is of great understanding: cipis 14:29 qui patiens est multa gubernatur prudentia qui

but he that is hasty of spirit exalteth folly.

14:30 A sound heart is the life of the flesh: but envy the autem inpatiens exaltat stultitiam suam 14:30 uita carnium

rottenness of the bones.

14:31 He that oppresseth the poor reproacheth his Maker: sanitas cordis putredo ossuum inuidia 14:31 qui calumnia-

but he that honoureth him hath mercy on the poor.

tur egentem exprobrat factori eius honorat autem eum qui

14:32 The wicked is driven away in his wickedness: but miseretur pauperis 14:32 in malitia sua expelletur impius

the righteous hath hope in his death.

14:33 Wisdom resteth in the heart of him that hath under- sperat autem iustus in morte sua 14:33 in corde prudentis

standing: but that which is in the midst of fools is made known.

requiescit sapientia et indoctos quoque erudiet 14:34 iusti-

14:34 Righteousness exalteth a nation: but sin is a reproach to any people.

14:35 The king’s favour is toward a wise servant: but his tia eleuat gentem miseros facit populos peccatum 14:35 ac-

wrath is against him that causeth shame.

ceptus est regi minister intellegens iracundiam eius inutilis sustinebit 15:1 responsio mollis frangit iram sermo durus suscitat fu- 15:1 A soft answer turneth away wrath: but grievous

words stir up anger.

rorem 15:2 lingua sapientium ornat scientiam os fatuorum 15:2 The tongue of the wise useth knowledge aright: but

the mouth of fools poureth out foolishness.

ebullit stultitiam 15:3 in omni loco oculi Domini contem- 15:3 The eyes of the LORD are in every place, beholding

the evil and the good.

plantur malos et bonos 15:4 lingua placabilis lignum uitae 15:4 A wholesome tongue is a tree of life: but perverse-

ness therein is a breach in the spirit.

quae inmoderata est conteret spiritum 15:5 stultus inridet 15:5 A fool despiseth his father’s instruction: but he that

regardeth reproof is prudent.

disciplinam patris sui qui autem custodit increpationes astutior fiet 15:6 domus iusti plurima fortitudo et in fructi- 15:6 In the house of the righteous is much treasure: but in

the revenues of the wicked is trouble.

bus impii conturbatur 15:7 labia sapientium disseminabunt 15:7 The lips of the wise disperse knowledge: but the heart

of the foolish doeth not so.

scientiam cor stultorum dissimile erit 15:8 uictimae impio- 15:8 The sacrifice of the wicked is an abomination to the

LORD: but the prayer of the upright is his delight.

rum abominabiles Domino uota iustorum placabilia 15:9 15:9 The way of the wicked is an abomination unto the

LORD: but he loveth him that followeth after righteousness.

abominatio est Domino uia impii qui sequitur iustitiam di-

ligetur ab eo 15:10 doctrina mala deserenti uiam qui in- 15:10 Correction is grievous unto him that forsaketh the

way: and he that hateth reproof shall die.

crepationes odit morietur 15:11 infernus et perditio coram 15:11 Hell and destruction are before the LORD: how

much more then the hearts of the children of men?

Domino quanto magis corda filiorum hominum 15:12 non 15:12 A scorner loveth not one that reproveth him: neither

will he go unto the wise.

amat pestilens eum qui se corripit nec ad sapientes graditur 15:13 cor gaudens exhilarat faciem in maerore animi 15:13 A merry heart maketh a cheerful countenance: but

by sorrow of the heart the spirit is broken.

deicitur spiritus 15:14 cor sapientis quaerit doctrinam et 15:14 The heart of him that hath understanding seeketh

knowledge: but the mouth of fools feedeth on foolishness.

os stultorum pascetur inperitia 15:15 omnes dies pauperis 15:15 All the days of the afflicted are evil: but he that is of

a merry heart hath a continual feast.

mali secura mens quasi iuge conuiuium 15:16 melius est 15:16 Better is little with the fear of the LORD than great

treasure and trouble therewith.

parum cum timore Domini quam thesauri magni et insatiabiles 15:17 melius est uocare ad holera cum caritate quam 15:17 Better is a dinner of herbs where love is, than a

stalled ox and hatred therewith.

ad uitulum saginatum cum odio 15:18 uir iracundus prouo- 15:18 A wrathful man stirreth up strife: but he that is slow

to anger appeaseth strife.

cat rixas qui patiens est mitigat suscitatas 15:19 iter pigro- 15:19 The way of the slothful man is as an hedge of thorns:

but the way of the righteous is made plain.

rum quasi sepes spinarum uia iustorum absque offendiculo 15:20 filius sapiens laetificat patrem et stultus homo despi- 15:20 A wise son maketh a glad father: but a foolish man

despiseth his mother.

cit matrem suam 15:21 stultitia gaudium stulto et uir pru- 15:21 Folly is joy to him that is destitute of wisdom: but a

man of understanding walketh uprightly.

dens dirigit gressus 15:22 dissipantur cogitationes ubi non 15:22 Without counsel purposes are disappointed: but in

the multitude of counsellors they are established.

est consilium ubi uero plures sunt consiliarii confirmantur 15:23 laetatur homo in sententia oris sui et sermo oportunus 15:23 A man hath joy by the answer of his mouth: and a

word spoken in due season, how good is it!

est optimus 15:24 semita uitae super eruditum ut declinet 15:24 The way of life is above to the wise, that he may

depart from hell beneath.

de inferno nouissimo 15:25 domum superborum demolietur 15:25 The LORD will destroy the house of the proud: but

he will establish the border of the widow.

Dominus et firmos facit terminos uiduae 15:26 abomina- 15:26 The thoughts of the wicked are an abomination to

the LORD: but the words of the pure are pleasant words.

tio Domini cogitationes malae et purus sermo pulcherrimus 15:27 He that is greedy of gain troubleth his own house; 15:27 conturbat domum suam qui sectatur auaritiam qui au-

but he that hateth gifts shall live.

15:28 The heart of the righteous studieth to answer: but tem odit munera uiuet 15:28 mens iusti meditatur oboedien-

the mouth of the wicked poureth out evil things.

15:29 The LORD is far from the wicked: but he heareth tiam os impiorum redundat malis 15:29 longe est Dominus

the prayer of the righteous.

15:30 The light of the eyes rejoiceth the heart: and a good ab impiis et orationes iustorum exaudiet 15:30 lux oculo-

report maketh the bones fat.

15:31 The ear that heareth the reproof of life abideth rum laetificat animam fama bona inpinguat ossa 15:31 auris

among the wise.

quae audit increpationes uitae in medio sapientium commo-

15:32 He that refuseth instruction despiseth his own soul: rabitur 15:32 qui abicit disciplinam despicit animam suam but he that heareth reproof getteth understanding.

15:33 The fear of the LORD is the instruction of wisdom; qui adquiescit increpationibus possessor est cordis 15:33 ti-

and before honour is humility.

mor Domini disciplina sapientiae et gloriam praecedit humilitas

16:1 The preparations of the heart in man, and the answer 16:1 hominis est animum praeparare et Dei gubernare lin-

of the tongue, is from the LORD.

16:2 All the ways of a man are clean in his own eyes; but guam 16:2 omnes uiae hominum patent oculis eius spiri-

the LORD weigheth the spirits.

16:3 Commit thy works unto the LORD, and thy thoughts tuum ponderator est Dominus 16:3 reuela Domino opera

shall be established.

16:4 The LORD hath made all things for himself: yea, tua et dirigentur cogitationes tuae 16:4 uniuersa propter se-

even the wicked for the day of evil.

met ipsum operatus est Dominus impium quoque ad diem

16:5 Every one that is proud in heart is an abomination malum 16:5 abominatio Domini omnis arrogans etiam si to the LORD: though hand join in hand, he shall not be unpunished.

manus ad manum fuerit non erit innocens 16:6 misericordia

16:6 By mercy and truth iniquity is purged: and by the fear of the LORD men depart from evil.

et ueritate rediiitur iniquitas et in timore Domini declinatur

16:7 When a man’s ways please the LORD, he maketh a malo 16:7 cum placuerint Domino uiae hominis inimicos

even his enemies to be at peace with him. 16:8 Better is a little with righteousness than great rev- quoque eius conuertet ad pacem 16:8 melius est parum cum enues without right.

16:9 A man’s heart deviseth his way: but the LORD di- iustitia quam multi fructus cum iniquitate 16:9 cor homi-

recteth his steps.

nis disponet uiam suam sed Domini est dirigere gressus eius

16:10 A divine sentence is in the lips of the king: his 16:10 diuinatio in labiis regis in iudicio non errabit os eius

mouth transgresseth not in judgment.

16:11 A just weight and balance are the LORD’s: all the 16:11 pondus et statera iudicia Domini sunt et opera eius om-

weights of the bag are his work.

16:12 It is an abomination to kings to commit wickedness: nes lapides sacculi 16:12 abominabiles regi qui agunt im-

for the throne is established by righteousness. 16:13 Righteous lips are the delight of kings; and they love pie quoniam iustitia firmatur solium 16:13 uoluntas regum him that speaketh right.

16:14 The wrath of a king is as messengers of death: but a labia iusta qui recta loquitur diligetur 16:14 indignatio re-

wise man will pacify it.

16:15 In the light of the king’s countenance is life; and his gis nuntii mortis et uir sapiens placabit eam 16:15 in hilari-

favour is as a cloud of the latter rain.

tate uultus regis uita et clementia eius quasi imber serotinus

16:16 How much better is it to get wisdom than gold! and 16:16 posside sapientiam quia auro melior est et adquire pru-

to get understanding rather to be chosen than silver!

16:17 The highway of the upright is to depart from evil: dentiam quia pretiosior est argento 16:17 semita iustorum

he that keepeth his way preserveth his soul. 16:18 Pride goeth before destruction, and an haughty spirit declinat mala custos animae suae seruat uiam suam 16:18 before a fall.

contritionem praecedit superbia et ante ruinam exaltatur spiritus 16:19 melius est humiliari cum mitibus quam diuidere 16:19 Better it is to be of an humble spirit with the lowly,

than to divide the spoil with the proud.

spolia cum superbis 16:20 eruditus in uerbo repperiet bona 16:20 He that handleth a matter wisely shall find good:

and whoso trusteth in the LORD, happy is he.

et qui in Domino sperat beatus est 16:21 qui sapiens corde 16:21 The wise in heart shall be called prudent: and the

sweetness of the lips increaseth learning.

est appellabitur prudens et qui dulcis eloquio maiora percipiet 16:22 fons uitae eruditio possidentis doctrina stultorum 16:22 Understanding is a wellspring of life unto him that

hath it: but the instruction of fools is folly.

fatuitas 16:23 cor sapientis erudiet os eius et labiis illius ad- 16:23 The heart of the wise teacheth his mouth, and addeth

learning to his lips.

det gratiam 16:24 fauus mellis uerba conposita dulcedo ani- 16:24 Pleasant words are as an honeycomb, sweet to the

soul, and health to the bones.

mae et sanitas ossuum 16:25 est uia quae uidetur homini re- 16:25 There is a way that seemeth right unto a man, but

the end thereof are the ways of death.

cta et nouissimum eius ducit ad mortem 16:26 anima labo- 16:26 He that laboureth laboureth for himself; for his

mouth craveth it of him.

rantis laborat sibi quia conpulit eum os suum 16:27 uir im- 16:27 An ungodly man diggeth up evil: and in his lips

there is as a burning fire.

pius fodit malum et in labiis eius ignis ardescit 16:28 homo 16:28 A froward man soweth strife: and a whisperer sepa-

rateth chief friends.

peruersus suscitat lites et uerbosus separat principes 16:29 16:29 A violent man enticeth his neighbour, and leadeth

him into the way that is not good.

uir iniquus lactat amicum suum et ducit eum per uiam non bonam 16:30 qui adtonitis oculis cogitat praua mordens la- 16:30 He shutteth his eyes to devise froward things: mov-

ing his lips he bringeth evil to pass.

bia sua perficit malum 16:31 corona dignitatis senectus in 16:31 The hoary head is a crown of glory, if it be found in

the way of righteousness.

uiis iustitiae repperietur 16:32 melior est patiens uiro forte 16:32 He that is slow to anger is better than the mighty;

and he that ruleth his spirit than he that taketh a city.

et qui dominatur animo suo expugnatore urbium 16:33 sor- 16:33 The lot is cast into the lap; but the whole disposing

thereof is of the LORD.

tes mittuntur in sinu sed a Domino temperantur

17:1 melior est buccella sicca cum gaudio quam domus 17:1 Better is a dry morsel, and quietness therewith, than

an house full of sacrifices with strife.

plena uictimis cum iurgio 17:2 seruus sapiens dominabi- 17:2 A wise servant shall have rule over a son that causeth

shame, and shall have part of the inheritance among the brethren.

tur filiis stultis et inter fratres hereditatem diuidet 17:3 sic-

17:3 The fining pot is for silver, and the furnace for gold: but the LORD trieth the hearts.

ut igne probatur argentum et aurum camino ita corda probat Dominus 17:4 malus oboedit linguae iniquae et fallax 17:4 A wicked doer giveth heed to false lips; and a liar

giveth ear to a naughty tongue.

obtemperat labiis mendacibus 17:5 qui despicit pauperem 17:5 Whoso mocketh the poor reproacheth his Maker: and

he that is glad at calamities shall not be unpunished.

exprobrat factori eius et qui in ruina laetatur alterius non erit inpunitus 17:6 corona senum filii filiorum et gloria fi- 17:6 Children’s children are the crown of old men; and the

glory of children are their fathers.

liorum patres sui 17:7 non decent stultum uerba conposita 17:7 Excellent speech becometh not a fool: much less do

lying lips a prince.

nec principem labium mentiens 17:8 gemma gratissima ex- 17:8 A gift is as a precious stone in the eyes of him that

hath it: whithersoever it turneth, it prospereth.

pectatio praestolantis quocumque se uerterit prudenter intellegit 17:9 qui celat delictum quaerit amicitias qui altero 17:9 He that covereth a transgression seeketh love; but he

that repeateth a matter separateth very friends.

sermone repetit separat foederatos 17:10 plus proficit cor- 17:10 A reproof entereth more into a wise man than an

hundred stripes into a fool.

reptio apud prudentem quam centum plagae apud stultum 17:11 semper iurgia quaerit malus angelus autem crudelis 17:11 An evil man seeketh only rebellion: therefore a

cruel messenger shall be sent against him.

17:12 Let a bear robbed of her whelps meet a man, rather mittetur contra eum 17:12 expedit magis ursae occurrere ra-

than a fool in his folly.

17:13 Whoso rewardeth evil for good, evil shall not depart ptis fetibus quam fatuo confidenti sibi in stultitia sua 17:13

from his house.

qui reddit mala pro bonis non recedet malum de domo eius

17:14 The beginning of strife is as when one letteth out 17:14 qui dimittit aquam caput est iurgiorum et antequam pa-

water: therefore leave off contention, before it be meddled with.

tiatur contumeliam iudicium deserit 17:15 et qui iustificat

17:15 He that justifieth the wicked, and he that condemneth the just, even they both are abomination to the LORD.

impium et qui condemnat iustum abominabilis est uterque

17:16 Wherefore is there a price in the hand of a fool to apud Dominum 17:16 quid prodeet habere diuitias stultum

get wisdom, seeing he hath no heart to it? 17:17 A friend loveth at all times, and a brother is born for cum sapientiam emere non possit 17:17 omni tempore diadversity.

17:18 A man void of understanding striketh hands, and ligit qui amicus est et frater in angustiis conprobatur 17:18

becometh surety in the presence of his friend.

homo stultus plaudet manibus cum spoponderit pro amico

17:19 He loveth transgression that loveth strife: and he suo 17:19 qui meditatur discordiam diligit rixas et qui ex-

that exalteth his gate seeketh destruction.

17:20 He that hath a froward heart findeth no good: and altat ostium quaerit ruinam 17:20 qui peruersi cordis est

he that hath a perverse tongue falleth into mischief.

non inueniet bonum et qui uertit linguam incidet in malum

17:21 He that begetteth a fool doeth it to his sorrow: and 17:21 natus est stultus in ignominiam suam sed nec pater in

the father of a fool hath no joy.

17:22 A merry heart doeth good like a medicine: but a fatuo laetabitur 17:22 animus gaudens aetatem floridam fa-

broken spirit drieth the bones.

17:23 A wicked man taketh a gift out of the bosom to per- cit spiritus tristis exsiccat ossa 17:23 munera de sinu impius

vert the ways of judgment.

17:24 Wisdom is before him that hath understanding; but accipit ut peruertat semitas iudicii 17:24 in facie prudentis

the eyes of a fool are in the ends of the earth.

17:25 A foolish son is a grief to his father, and bitterness lucet sapientia oculi stultorum in finibus terrae 17:25 ira pa-

to her that bare him.

17:26 Also to punish the just is not good, nor to strike tris filius stultus et dolor matris quae genuit eum 17:26 non

princes for equity.

est bonum damnum inferre iusto nec percutere principem

17:27 He that hath knowledge spareth his words: and a qui recta iudicat 17:27 qui moderatur sermones suos doctus

man of understanding is of an excellent spirit.

17:28 Even a fool, when he holdeth his peace, is counted et prudens est et pretiosi spiritus uir eruditus 17:28 stultus

wise: and he that shutteth his lips is esteemed a man of understanding.

quoque si tacuerit sapiens putabitur et si conpresserit labia sua intellegens

18:1 Through desire a man, having separated himself, 18:1 occasiones quaerit qui uult recedere ab amico omni

seeketh and intermeddleth with all wisdom.

18:2 A fool hath no delight in understanding, but that his tempore erit exprobrabilis 18:2 non recipit stultus uerba

heart may discover itself.

18:3 When the wicked cometh, then cometh also con- prudentiae nisi ea dixeris quae uersantur in corde eius 18:3

tempt, and with ignominy reproach.

impius cum in profundum uenerit peccatorum contemnit sed

18:4 The words of a man’s mouth are as deep waters, and sequitur eum ignominia et obprobrium 18:4 aqua profunda

the wellspring of wisdom as a flowing brook.

18:5 It is not good to accept the person of the wicked, to uerba ex ore uiri et torrens redundans fons sapientiae 18:5

overthrow the righteous in judgment.

accipere personam impii non est bonum ut declines a ueritate

18:6 A fool’s lips enter into contention, and his mouth cal- iudicii 18:6 labia stulti inmiscunt se rixis et os eius iurgia

leth for strokes.

18:7 A fool’s mouth is his destruction, and his lips are the prouocat 18:7 os stulti contritio eius et labia illius ruina ani-

snare of his soul.

mae eius 18:8 uerba bilinguis quasi simplicia et ipsa perue- 18:8 The words of a talebearer are as wounds, and they go

down into the innermost parts of the belly.

niunt usque ad interiora uentris 18:9 qui mollis et dissolutus 18:9 He also that is slothful in his work is brother to him

that is a great waster.

est in opere suo frater est sua opera dissipantis 18:10 turris 18:10 The name of the LORD is a strong tower: the righ-

teous runneth into it, and is safe.

fortissima nomen Domini ad ipsum currit iustus et exaltabitur 18:11 substantia diuitis urbs roboris eius et quasi murus 18:11 The rich man’s wealth is his strong city, and as an

high wall in his own conceit.

ualidus circumdans eum 18:12 antequam conteratur exalta- 18:12 Before destruction the heart of man is haughty, and

before honour is humility.

tur cor hominis et antequam glorificetur humiliatur 18:13 18:13 He that answereth a matter before he heareth it, it is

folly and shame unto him.

qui prius respondit quam audiat stultum se esse demonstrat et confusione dignum 18:14 spiritus uiri sustentat inbecilli- 18:14 The spirit of a man will sustain his infirmity; but a

wounded spirit who can bear?

tatem suam spiritum uero ad irascendum facilem quis poterit sustinere 18:15 cor prudens possidebit scientiam et auris 18:15 The heart of the prudent getteth knowledge; and the

ear of the wise seeketh knowledge.

sapientium quaerit doctrinam 18:16 donum hominis dila- 18:16 A man’s gift maketh room for him, and bringeth him

before great men.

tat uiam eius et ante principes spatium ei facit 18:17 iustus 18:17 He that is first in his own cause seemeth just; but his

neighbour cometh and searcheth him.

prior est accusator sui uenit amicus eius et inuestigauit eum 18:18 contradictiones conprimit sors et inter potentes quo- 18:18 The lot causeth contentions to cease, and parteth

between the mighty.

que diiudicat 18:19 frater qui adiuuatur a fratre quasi ciuitas 18:19 A brother offended is harder to be won than a strong

city: and their contentions are like the bars of a castle.

firma et iudicia quasi uectes urbium 18:20 de fructu oris uiri 18:20 A man’s belly shall be satisfied with the fruit of his

mouth; and with the increase of his lips shall he be filled.

replebitur uenter eius et genimina labiorum illius saturabunt eum 18:21 mors et uita in manu linguae qui diligunt eam 18:21 Death and life are in the power of the tongue: and

they that love it shall eat the fruit thereof.

comedent fructus eius 18:22 qui inuenit mulierem inuenit 18:22 Whoso findeth a wife findeth a good thing, and ob-

taineth favour of the LORD.

bonum et hauriet iucunditatem a Domino 18:23 cum obse- 18:23 The poor useth intreaties; but the rich answereth

roughly.

crationibus loquetur pauper et diues effabitur rigide 18:24 18:24 A man that hath friends must shew himself friendly:

and there is a friend that sticketh closer than a brother.

uir amicalis ad societatem magis amicus erit quam frater

19:1 melior est pauper qui ambulat in simplicitate sua 19:1 Better is the poor that walketh in his integrity, than

he that is perverse in his lips, and is a fool.

quam torquens labia insipiens 19:2 ubi non est scientia ani- 19:2 Also, that the soul be without knowledge, it is not

good; and he that hasteth with his feet sinneth.

mae non est bonum et qui festinus est pedibus offendit 19:3 19:3 The foolishness of man perverteth his way: and his

heart fretteth against the LORD.

stultitia hominis subplantat gressus eius et contra Deum feruet animo suo 19:4 diuitiae addunt amicos plurimos a pau- 19:4 Wealth maketh many friends; but the poor is sepa-

rated from his neighbour.

pere autem et hii quos habuit separantur 19:5 testis fal- 19:5 A false witness shall not be unpunished, and he that

speaketh lies shall not escape.

sus non erit inpunitus et qui mendacia loquitur non effugiet 19:6 multi colunt personam potentis et amici sunt dona tri- 19:6 Many will intreat the favour of the prince: and every

man is a friend to him that giveth gifts.

buenti 19:7 fratres hominis pauperis oderunt eum insuper 19:7 All the brethren of the poor do hate him: how much

more do his friends go far from him? he pursueth them with words, yet they are wanting to him.

et amici procul recesserunt ab eo qui tantum uerba secta-19:8 He that getteth wisdom loveth his own soul: he that tur nihil habebit 19:8 qui autem possessor est mentis diligit

keepeth understanding shall find good.

19:9 A false witness shall not be unpunished, and he that animam suam et custos prudentiae inueniet bona 19:9 tes-

speaketh lies shall perish.

tis falsus non erit inpunitus et qui loquitur mendacia peribit

19:10 Delight is not seemly for a fool; much less for a 19:10 non decent stultum deliciae nec seruum dominari prin-

servant to have rule over princes.

19:11 The discretion of a man deferreth his anger; and it cipibus 19:11 doctrina uiri per patientiam noscitur et gloria

is his glory to pass over a transgression.

19:12 The king’s wrath is as the roaring of a lion; but his eius est iniqua praetergredi 19:12 sicut fremitus leonis ita

favour is as dew upon the grass.

19:13 A foolish son is the calamity of his father: and the et regis ira et sicut ros super herbam ita hilaritas eius 19:13

contentions of a wife are a continual dropping.

dolor patris filius stultus et tecta iugiter perstillantia litigiosa

19:14 House and riches are the inheritance of fathers: and mulier 19:14 domus et diuitiae dantur a patribus a Domino a prudent wife is from the LORD.

19:15 Slothfulness casteth into a deep sleep; and an idle autem proprie uxor prudens 19:15 pigredo inmittit soporem

soul shall suffer hunger.

19:16 He that keepeth the commandment keepeth his own et anima dissoluta esuriet 19:16 qui custodit mandatum cu-

soul; but he that despiseth his ways shall die.

stodit animam suam qui autem neglegit uias suas mortifica-

19:17 He that hath pity upon the poor lendeth unto the bitur 19:17 feneratur Domino qui miseretur pauperis et ui-

LORD; and that which he hath given will he pay him again.

cissitudinem suam reddet ei 19:18 erudi filium tuum ne de-

19:18 Chasten thy son while there is hope, and let not thy soul spare for his crying.

speres ad interfectionem autem eius ne ponas animam tuam

19:19 A man of great wrath shall suffer punishment: for if 19:19 qui inpatiens est sustinebit damnum et cum rapuerit al-

thou deliver him, yet thou must do it again.

19:20 Hear counsel, and receive instruction, that thou iud adponet 19:20 audi consilium et suscipe disciplinam ut

mayest be wise in thy latter end.

19:21 There are many devices in a man’s heart; neverthe- sis sapiens in nouissimis tuis 19:21 multae cogitationes in

less the counsel of the LORD, that shall stand.

19:22 The desire of a man is his kindness: and a poor man corde uiri uoluntas autem Domini permanebit 19:22 homo

is better than a liar.

indigens misericors est et melior pauper quam uir mendax

19:23 The fear of the LORD tendeth to life: and he that 19:23 timor Domini ad uitam et in plenitudine commorabitur

hath it shall abide satisfied; he shall not be visited with evil.

absque uisitatione pessimi 19:24 abscondit piger manum

19:24 A slothful man hideth his hand in his bosom, and will not so much as bring it to his mouth again. 19:25 Smite a scorner, and the simple will beware: and re- suam sub ascella nec ad os suum adplicat eam 19:25 pesprove one that hath understanding, and he will understand knowledge.

tilente flagellato stultus sapientior erit sin autem corripue-

19:26 He that wasteth his father, and chaseth away his ris sapientem intelleget disciplinam 19:26 qui adfligit pat-

mother, is a son that causeth shame, and bringeth reproach. 19:27 Cease, my son, to hear the instruction that causeth rem et fugat matrem ignominiosus est et infelix 19:27 non to err from the words of knowledge.

cesses fili audire doctrinam nec ignores sermones scientiae

19:28 An ungodly witness scorneth judgment: and the 19:28 testis iniquus deridet iudicium et os impiorum deuorat

mouth of the wicked devoureth iniquity.

19:29 Judgments are prepared for scorners, and stripes for iniquitatem 19:29 parata sunt derisoribus iudicia et mallei

the back of fools.

percutientes stultorum corporibus

20:1 Wine is a mocker, strong drink is raging: and whoso- 20:1 luxuriosa res uinum et tumultuosa ebrietas quicum-

ever is deceived thereby is not wise.

20:2 The fear of a king is as the roaring of a lion: whoso que his delectatur non erit sapiens 20:2 sicut rugitus leo-

provoketh him to anger sinneth against his own soul.

nis ita terror regis qui prouocat eum peccat in animam suam 20:3 honor est homini qui separat se a contentionibus omnes 20:3 It is an honour for a man to cease from strife: but

every fool will be meddling.

autem stulti miscentur contumeliis 20:4 propter frigus piger 20:4 The sluggard will not plow by reason of the cold;

therefore shall he beg in harvest, and have nothing.

arare noluit mendicabit ergo aestate et non dabitur ei 20:5 20:5 Counsel in the heart of man is like deep water; but a

man of understanding will draw it out.

sicut aqua profunda sic consilium in corde uiri sed homo sapiens exhauriet illud 20:6 multi homines misericordes uo- 20:6 Most men will proclaim every one his own goodness:

but a faithful man who can find?

cantur uirum autem fidelem quis inueniet 20:7 iustus qui 20:7 The just man walketh in his integrity: his children are

blessed after him.

ambulat in simplicitate sua beatos post se filios derelinquet

20:8 rex qui sedet in solio iudicii dissipat omne malum in- 20:8 A king that sitteth in the throne of judgment scattereth

away all evil with his eyes.

tuitu suo 20:9 quis potest dicere mundum est cor meum pu- 20:9 Who can say, I have made my heart clean, I am pure

from my sin?

rus sum a peccato 20:10 pondus et pondus mensura et men- 20:10 Divers weights, and divers measures, both of them

are alike abomination to the LORD.

sura utrumque abominabile est apud Deum 20:11 ex stu- 20:11 Even a child is known by his doings, whether his

work be pure, and whether it be right.

diis suis intellegitur puer si munda et si recta sint opera eius 20:12 aurem audientem et oculum uidentem Dominus fecit 20:12 The hearing ear, and the seeing eye, the LORD hath

made even both of them.

utrumque 20:13 noli diligere somnum ne te egestas oppri- 20:13 Love not sleep, lest thou come to poverty; open

thine eyes, and thou shalt be satisfied with bread.

mat aperi oculos tuos et saturare panibus 20:14 malum est 20:14 It is naught, it is naught, saith the buyer: but when

he is gone his way, then he boasteth.

malum est dicit omnis emptor et cum recesserit tunc glo-

riabitur 20:15 est aurum et multitudo gemmarum uas au- 20:15 There is gold, and a multitude of rubies: but the lips

of knowledge are a precious jewel.

tem pretiosum labia scientiae 20:16 tolle uestimentum eius 20:16 Take his garment that is surety for a stranger: and

take a pledge of him for a strange woman.

qui fideiussor extitit alieni et pro extraneis aufer pignus ab

eo 20:17 suauis est homini panis mendacii et postea imp- 20:17 Bread of deceit is sweet to a man; but afterwards his

mouth shall be filled with gravel.

lebitur os eius calculo 20:18 cogitationes consiliis roboran- 20:18 Every purpose is established by counsel: and with

good advice make war.

tur et gubernaculis tractanda sunt bella 20:19 ei qui reue- 20:19 He that goeth about as a talebearer revealeth secrets:

therefore meddle not with him that flattereth with his lips.

lat mysteria et ambulat fraudulenter et dilatat labia sua ne commiscearis 20:20 qui maledicit patri suo et matri extin- 20:20 Whoso curseth his father or his mother, his lamp

shall be put out in obscure darkness.

guetur lucerna eius in mediis tenebris 20:21 hereditas ad 20:21 An inheritance may be gotten hastily at the begin-

ning; but the end thereof shall not be blessed.

quam festinatur in principio in nouissimo benedictione care-

bit 20:22 ne dicas reddam malum expecta Dominum et libe- 20:22 Say not thou, I will recompense evil; but wait on the

LORD, and he shall save thee.

rabit te 20:23 abominatio est apud Deum pondus et pondus 20:23 Divers weights are an abomination unto the LORD;

and a false balance is not good.

statera dolosa non est bona 20:24 a Domino diriguntur gres- 20:24 Man’s goings are of the LORD; how can a man then

understand his own way?

sus uiri quis autem hominum intellegere potest uiam suam 20:25 ruina est hominis deuorare sanctos et post uota trac- 20:25 It is a snare to the man who devoureth that which is

holy, and after vows to make enquiry.

tare 20:26 dissipat impios rex sapiens et curuat super eos 20:26 A wise king scattereth the wicked, and bringeth the

wheel over them.

fornicem 20:27 lucerna Domini spiraculum hominis quae 20:27 The spirit of man is the candle of the LORD, search-

ing all the inward parts of the belly.

20:28 Mercy and truth preserve the king: and his throne is inuestigat omnia secreta uentris 20:28 misericordia et ue-

upholden by mercy.

ritas custodiunt regem et roboratur clementia thronus eius

20:29 The glory of young men is their strength: and the 20:29 exultatio iuuenum fortitudo eorum et dignitas senum

beauty of old men is the grey head.

20:30 The blueness of a wound cleanseth away evil: so do canities 20:30 liuor uulneris absterget mala et plagae in se-

stripes the inward parts of the belly.

cretioribus uentris

21:1 The king’s heart is in the hand of the LORD, as the 21:1 sicut diuisiones aquarum ita cor regis in manu Domini

rivers of water: he turneth it whithersoever he will.

21:2 Every way of a man is right in his own eyes: but the quocumque uoluerit inclinabit illud 21:2 omnis uia uiri re-

LORD pondereth the hearts.

21:3 To do justice and judgment is more acceptable to the cta sibi uidetur adpendit autem corda Dominus 21:3 facere

LORD than sacrifice.

misericordiam et iudicium magis placent Domino quam uic-

21:4 An high look, and a proud heart, and the plowing of timae 21:4 exaltatio oculorum et dilatatio cordis lucerna

the wicked, is sin. 21:5 The thoughts of the diligent tend only to plenteous- impiorum peccatum 21:5 cogitationes robusti semper in ness; but of every one that is hasty only to want. 21:6 The getting of treasures by a lying tongue is a vanity abundantia omnis autem piger semper in egestate 21:6 qui tossed to and fro of them that seek death.

congregat thesauros lingua mendacii uanus est et inpinge-

21:7 The robbery of the wicked shall destroy them; be- tur ad laqueos mortis 21:7 rapinae impiorum detrahent eos

cause they refuse to do judgment.

21:8 The way of man is froward and strange: but as for the quia noluerunt facere iudicium 21:8 peruersa uia uiri aliena

pure, his work is right. 21:9 It is better to dwell in a corner of the housetop, than est qui autem mundus est rectum opus eius 21:9 melius est with a brawling woman in a wide house.

sedere in angulo domatis quam cum muliere litigiosa et in

21:10 The soul of the wicked desireth evil: his neighbour domo communi 21:10 anima impii desiderat malum non findeth no favour in his eyes.

21:11 When the scorner is punished, the simple is made miserebitur proximo suo 21:11 multato pestilente sapien-

wise: and when the wise is instructed, he receiveth knowledge.

tior erit paruulus et si sectetur sapientem sumet scientiam

21:12 The righteous man wisely considereth the house of 21:12 excogitat iustus de domo impii ut detrahat impios in

the wicked: but God overthroweth the wicked for their wickedness.

malum 21:13 qui obturat aurem suam ad clamorem paupe-

21:13 Whoso stoppeth his ears at the cry of the poor, he also shall cry himself, but shall not be heard.

21:14 A gift in secret pacifieth anger: and a reward in the ris et ipse clamabit et non exaudietur 21:14 munus abscon-

bosom strong wrath.

ditum extinguet iras et donum in sinu indignationem maxi-

21:15 It is joy to the just to do judgment: but destruction mam 21:15 gaudium iusto est facere iudicium et pauor opeshall be to the workers of iniquity.

21:16 The man that wandereth out of the way of under- rantibus iniquitatem 21:16 uir qui errauerit a uia doctrinae

standing shall remain in the congregation of the dead.

21:17 He that loveth pleasure shall be a poor man: he that in coetu gigantum commorabitur 21:17 qui diligit epulas in

loveth wine and oil shall not be rich.

21:18 The wicked shall be a ransom for the righteous, and egestate erit qui amat uinum et pinguia non ditabitur 21:18

the transgressor for the upright.

21:19 It is better to dwell in the wilderness, than with a pro iusto datur impius et pro rectis iniquus 21:19 melius

contentious and an angry woman.

est habitare in terra deserta quam cum muliere rixosa et ira-

21:20 There is treasure to be desired and oil in the cunda 21:20 thesaurus desiderabilis et oleum in habitaculo

dwelling of the wise; but a foolish man spendeth it up.

21:21 He that followeth after righteousness and mercy iusti et inprudens homo dissipabit illud 21:21 qui sequitur

findeth life, righteousness, and honour.

iustitiam et misericordiam inueniet uitam et iustitiam et gloriam 21:22 ciuitatem fortium ascendit sapiens et destruxit 21:22 A wise man scaleth the city of the mighty, and

casteth down the strength of the confidence thereof.

robur fiduciae eius 21:23 qui custodit os suum et linguam 21:23 Whoso keepeth his mouth and his tongue keepeth

his soul from troubles.

suam custodit ab angustiis animam suam 21:24 superbus 21:24 Proud and haughty scorner is his name, who dealeth

in proud wrath.

et arrogans uocatur indoctus qui in ira operatur superbiam 21:25 desideria occidunt pigrum noluerunt enim quicquam 21:25 The desire of the slothful killeth him; for his hands

refuse to labour.

manus eius operari 21:26 tota die concupiscit et desiderat 21:26 He coveteth greedily all the day long: but the righ-

teous giveth and spareth not.

qui autem iustus est tribuet et non cessabit 21:27 hostiae 21:27 The sacrifice of the wicked is abomination: how

much more, when he bringeth it with a wicked mind?

impiorum abominabiles quia offeruntur ex scelere 21:28 te- 21:28 A false witness shall perish: but the man that heareth

speaketh constantly.

stis mendax peribit uir oboediens loquitur uictoriam 21:29 21:29 A wicked man hardeneth his face: but as for the

upright, he directeth his way.

uir impius procaciter obfirmat uultum suum qui autem rectus est corrigit uiam suam 21:30 non est sapientia non est 21:30 There is no wisdom nor understanding nor counsel

against the LORD.

prudentia non est consilium contra Dominum 21:31 equus 21:31 The horse is prepared against the day of battle: but

safety is of the LORD.

paratur ad diem belli Dominus autem salutem tribuet

22:1 melius est nomen bonum quam diuitiae multae super 22:1 A GOOD name is rather to be chosen than great

riches, and loving favour rather than silver and gold.

argentum et aurum gratia bona 22:2 diues et pauper obuia- 22:2 The rich and poor meet together: the LORD is the

maker of them all.

uerunt sibi utriusque operator est Dominus 22:3 callidus ui- 22:3 A prudent man foreseeth the evil, and hideth himself:

but the simple pass on, and are punished.

dit malum et abscondit se innocens pertransiit et adflictus est damno 22:4 finis modestiae timor Domini diuitiae et glo- 22:4 By humility and the fear of the LORD are riches, and

honour, and life.

ria et uita 22:5 arma et gladii in uia peruersi custos animae 22:5 Thorns and snares are in the way of the froward: he

that doth keep his soul shall be far from them.

suae longe recedit ab eis 22:6 prouerbium est adulescens 22:6 Train up a child in the way he should go: and when

he is old, he will not depart from it.

iuxta uiam suam etiam cum senuerit non recedet ab ea 22:7 22:7 The rich ruleth over the poor, and the borrower is

servant to the lender.

diues pauperibus imperat et qui accipit mutuum seruus est fenerantis 22:8 qui seminat iniquitatem metet mala et uirga 22:8 He that soweth iniquity shall reap vanity: and the rod

of his anger shall fail.

irae suae consummabitur 22:9 qui pronus est ad misericor- 22:9 He that hath a bountiful eye shall be blessed; for he

giveth of his bread to the poor.

diam benedicetur de panibus enim suis dedit pauperi 22:10 22:10 Cast out the scorner, and contention shall go out;

yea, strife and reproach shall cease.

eice derisorem et exibit cum eo iurgium cessabuntque causae et contumeliae 22:11 qui diligit cordis munditiam prop- 22:11 He that loveth pureness of heart, for the grace of his

lips the king shall be his friend.

ter gratiam labiorum suorum habebit amicum regem 22:12 22:12 The eyes of the LORD preserve knowledge, and he

overthroweth the words of the transgressor.

oculi Domini custodiunt scientiam et subplantantur uerba iniqui 22:13 dicit piger leo foris in medio platearum occiden- 22:13 The slothful man saith, There is a lion without, I

shall be slain in the streets.

dus sum 22:14 fouea profunda os alienae cui iratus est Do- 22:14 The mouth of strange women is a deep pit: he that

is abhorred of the LORD shall fall therein.

minus incidet in eam 22:15 stultitia conligata est in corde 22:15 Foolishness is bound in the heart of a child; but the

rod of correction shall drive it far from him.

pueri et uirga disciplinae fugabit eam 22:16 qui calumniatur 22:16 He that oppresseth the poor to increase his riches,

and he that giveth to the rich, shall surely come to want.

pauperem ut augeat diuitias suas dabit ipse ditiori et egebit

22:17 Bow down thine ear, and hear the words of the wise, 22:17 inclina aurem tuam et audi uerba sapientium adpone

and apply thine heart unto my knowledge.

22:18 For it is a pleasant thing if thou keep them within autem cor ad doctrinam meam 22:18 quae pulchra erit tibi

thee; they shall withal be fitted in thy lips.

cum seruaueris eam in uentre tuo et redundabit in labiis tuis

22:19 That thy trust may be in the LORD, I have made 22:19 ut sit in Domino fiducia tua unde et ostendi eam tibi ho-

known to thee this day, even to thee.

22:20 Have not I written to thee excellent things in coun- die 22:20 ecce descripsi eam tibi tripliciter in cogitationibus

sels and knowledge,

22:21 That I might make thee know the certainty of the et scientia 22:21 ut ostenderem tibi firmitatem et eloquia

words of truth; that thou mightest answer the words of truth to them that send unto thee?

ueritatis respondere ex his illi qui misit te 22:22 non facias

22:22 Rob not the poor, because he is poor: neither oppress the afflicted in the gate:

uiolentiam pauperi quia pauper est neque conteras egenum

22:23 For the LORD will plead their cause, and spoil the in porta 22:23 quia Dominus iudicabit causam eius et confi-

soul of those that spoiled them.

22:24 Make no friendship with an angry man; and with a get eos qui confixerint animam eius 22:24 noli esse amicus

furious man thou shalt not go: 22:25 Lest thou learn his ways, and get a snare to thy soul. homini iracundo neque ambules cum uiro furioso 22:25 ne

forte discas semitas eius et sumas scandalum animae tuae

22:26 Be not thou one of them that strike hands, or of them 22:26 noli esse cum his qui defigunt manus suas et qui uathat are sureties for debts. 22:27 If thou hast nothing to pay, why should he take away des se offerunt pro debitis 22:27 si enim non habes unde thy bed from under thee?

restituas quid causae est ut tollat operimentum de cubili tuo

22:28 Remove not the ancient landmark, which thy fathers 22:28 ne transgrediaris terminos antiquos quos posuerunt pa-

have set.

22:29 Seest thou a man diligent in his business? he shall tres tui 22:29 uidisti uirum uelocem in opere suo coram re-

stand before kings; he shall not stand before mean men.

gibus stabit nec erit ante ignobiles

23:1 When thou sittest to eat with a ruler, consider dili- 23:1 quando sederis ut comedas cum principe diligenter gently what is before thee:

23:2 And put a knife to thy throat, if thou be a man given adtende quae posita sunt ante faciem tuam 23:2 et statue

to appetite.

cultrum in gutture tuo si tamen habes in potestate animam

23:3 Be not desirous of his dainties: for they are deceitful tuam 23:3 ne desideres de cibis eius in quo est panis menmeat.

23:4 Labour not to be rich: cease from thine own wisdom. dacii 23:4 noli laborare ut diteris sed prudentiae tuae pone

23:5 Wilt thou set thine eyes upon that which is not? for modum 23:5 ne erigas oculos tuos ad opes quas habere riches certainly make themselves wings; they fly away as an eagle toward heaven.

non potes quia facient sibi pinnas quasi aquilae et auola-

23:6 Eat thou not the bread of him that hath an evil eye, bunt in caelum 23:6 ne comedas cum homine inuido et ne neither desire thou his dainty meats:

23:7 For as he thinketh in his heart, so is he: Eat and drink, desideres cibos eius 23:7 quoniam in similitudinem arioli

saith he to thee; but his heart is not with thee.

et coniectoris aestimat quod ignorat comede et bibe dicet

23:8 The morsel which thou hast eaten shalt thou vomit tibi et mens eius non est tecum 23:8 cibos quos comederas

up, and lose thy sweet words. 23:9 Speak not in the ears of a fool: for he will despise the euomes et perdes pulchros sermones tuos 23:9 in auribus wisdom of thy words.

insipientium ne loquaris quia despicient doctrinam eloquii

23:10 Remove not the old landmark; and enter not into the tui 23:10 ne adtingas terminos paruulorum et agrum pupil-

fields of the fatherless:

lorum ne introeas 23:11 propinquus enim eorum Fortis est 23:11 For their redeemer is mighty; he shall plead their

cause with thee.

et ipse iudicabit contra te causam illorum 23:12 ingredia- 23:12 Apply thine heart unto instruction, and thine ears to

the words of knowledge.

tur ad doctrinam cor tuum et aures tuae ad uerba scientiae 23:13 noli subtrahere a puero disciplinam si enim percusse- 23:13 Withhold not correction from the child: for if thou

beatest him with the rod, he shall not die.

ris eum uirga non morietur 23:14 tu uirga percuties eum et 23:14 Thou shalt beat him with the rod, and shalt deliver

his soul from hell.

animam eius de inferno liberabis 23:15 fili mi si sapiens 23:15 My son, if thine heart be wise, my heart shall re-

joice, even mine.

fuerit animus tuus gaudebit tecum cor meum 23:16 et exul- 23:16 Yea, my reins shall rejoice, when thy lips speak right

things.

tabunt renes mei cum locuta fuerint rectum labia tua 23:17 23:17 Let not thine heart envy sinners: but be thou in the

fear of the LORD all the day long.

non aemuletur cor tuum peccatores sed in timore Domini esto tota die 23:18 quia habebis spem in nouissimo et prae- 23:18 For surely there is an end; and thine expectation

shall not be cut off.

stolatio tua non auferetur 23:19 audi fili mi et esto sapiens 23:19 Hear thou, my son, and be wise, and guide thine

heart in the way.

et dirige in uia animum tuum 23:20 noli esse in conuiuiis 23:20 Be not among winebibbers; among riotous eaters of

flesh:

potatorum nec in comesationibus eorum qui carnes ad uescendum conferunt 23:21 quia uacantes potibus et dantes 23:21 For the drunkard and the glutton shall come to

poverty: and drowsiness shall clothe a man with rags.

symbola consumentur et uestietur pannis dormitatio 23:22 23:22 Hearken unto thy father that begat thee, and despise

not thy mother when she is old.

audi patrem tuum qui genuit te et ne contemnas cum senuerit mater tua 23:23 ueritatem eme et noli uendere sapientiam et 23:23 Buy the truth, and sell it not; also wisdom, and in-

struction, and understanding.

doctrinam et intellegentiam 23:24 exultat gaudio pater iusti 23:24 The father of the righteous shall greatly rejoice: and

he that begetteth a wise child shall have joy of him.

qui sapientem genuit laetabitur in eo 23:25 gaudeat pater 23:25 Thy father and thy mother shall be glad, and she that

bare thee shall rejoice.

tuus et mater tua et exultet quae genuit te 23:26 praebe fili 23:26 My son, give me thine heart, and let thine eyes ob-

serve my ways.

mi cor tuum mihi et oculi tui uias meas custodiant 23:27 23:27 For a whore is a deep ditch; and a strange woman is

a narrow pit.

fouea enim profunda est meretrix et puteus angustus aliena 23:28 insidiatur in uia quasi latro et quos incautos uiderit in- 23:28 She also lieth in wait as for a prey, and increaseth

the transgressors among men.

terficit 23:29 cui uae cuius patri uae cui rixae cui foueae cui 23:29 Who hath woe? who hath sorrow? who hath con-

tentions? who hath babbling? who hath wounds without cause? who hath redness of eyes?

sine causa uulnera cui suffusio oculorum 23:30 nonne his

23:30 They that tarry long at the wine; they that go to seek mixed wine.

qui morantur in uino et student calicibus epotandis 23:31 23:31 Look not thou upon the wine when it is red, when it

giveth his colour in the cup, when it moveth itself aright.

ne intuearis uinum quando flauescit cum splenduerit in uitro color eius ingreditur blande 23:32 sed in nouissimo morde- 23:32 At the last it biteth like a serpent, and stingeth like

an adder.

bit ut coluber et sicut regulus uenena diffundet 23:33 oculi 23:33 Thine eyes shall behold strange women, and thine

heart shall utter perverse things.

tui uidebunt extraneas et cor tuum loquetur peruersa 23:34 23:34 Yea, thou shalt be as he that lieth down in the midst

of the sea, or as he that lieth upon the top of a mast.

et eris sicut dormiens in medio mari et quasi sopitus gubernator amisso clauo 23:35 et dices uerberauerunt me sed non 23:35 They have stricken me, shalt thou say, and I was not

sick; they have beaten me, and I felt it not: when shall I awake? I will seek it yet again.

dolui traxerunt me et ego non sensi quando euigilabo et rur-

sum uina repperiam

24:1 Be not thou envious against evil men, neither desire 24:1 ne aemuleris uiros malos nec desideres esse cum eis to be with them. 24:2 For their heart studieth destruction, and their lips talk 24:2 quia rapinas meditatur mens eorum et fraudes labia eoof mischief.

24:3 Through wisdom is an house builded; and by under- rum loquuntur 24:3 sapientia aedificabitur domus et pru-

standing it is established:

24:4 And by knowledge shall the chambers be filled with dentia roborabitur 24:4 in doctrina replebuntur cellaria uni-

all precious and pleasant riches.

24:5 A wise man is strong; yea, a man of knowledge in- uersa substantia pretiosa et pulcherrima 24:5 uir sapiens et

creaseth strength.

24:6 For by wise counsel thou shalt make thy war: and in fortis est et uir doctus robustus et ualidus 24:6 quia cum

multitude of counsellors there is safety.

dispositione initur bellum et erit salus ubi multa consilia

24:7 Wisdom is too high for a fool: he openeth not his sunt 24:7 excelsa stulto sapientia in porta non aperiet os

mouth in the gate.

24:8 He that deviseth to do evil shall be called a mis- suum 24:8 qui cogitat malefacere stultus uocabitur 24:9

chievous person. 24:9 The thought of foolishness is sin: and the scorner is an abomination to men.

cogitatio stulti peccatum est et abominatio hominum detrac-

24:10 If thou faint in the day of adversity, thy strength is tor 24:10 si desperaueris lassus in die angustiae inminuetur

small.

24:11 If thou forbear to deliver them that are drawn unto fortitudo tua 24:11 erue eos qui ducuntur ad mortem et qui

death, and those that are ready to be slain;

24:12 If thou sayest, Behold, we knew it not; doth not he trahuntur ad interitum liberare ne cesses 24:12 si dixeris ui-

that pondereth the heart consider it? and he that keepeth thy soul, doth not he know it? and shall not he render to

every man according to his works? res non suppetunt qui inspector est cordis ipse intellegit et seruatorem animae tuae nihil fallit reddetque homini iuxta

24:13 My son, eat thou honey, because it is good; and the opera sua 24:13 comede fili mi mel quia bonum est et fahoneycomb, which is sweet to thy taste:

24:14 So shall the knowledge of wisdom be unto thy soul: uum dulcissimum gutturi tuo 24:14 sic et doctrina sapien-

when thou hast found it, then there shall be a reward, and thy expectation shall not be cut off.

tiae animae tuae quam cum inueneris habebis in nouissimis

24:15 Lay not wait, O wicked man, against the dwelling et spes tua non peribit 24:15 ne insidieris et quaeras impie-

of the righteous; spoil not his resting place:

24:16 For a just man falleth seven times, and riseth up tatem in domo iusti neque uastes requiem eius 24:16 septies

again: but the wicked shall fall into mischief.

enim cadet iustus et resurget impii autem corruent in malum

24:17 Rejoice not when thine enemy falleth, and let not 24:17 cum ceciderit inimicus tuus ne gaudeas et in ruina eius

thine heart be glad when he stumbleth:

24:18 Lest the LORD see it, and it displease him, and he ne exultet cor tuum 24:18 ne forte uideat Dominus et dis-

turn away his wrath from him. 24:19 Fret not thyself because of evil men, neither be thou pliceat ei et auferat ab eo iram suam 24:19 ne contendas envious at the wicked:

24:20 For there shall be no reward to the evil man; the cum pessimis nec aemuleris impios 24:20 quoniam non ha-

candle of the wicked shall be put out.

bent futurorum spem mali et lucerna impiorum extinguetur

24:21 My son, fear thou the LORD and the king: and med- 24:21 time Dominum fili mi et regem et cum detractoribus

dle not with them that are given to change:

24:22 For their calamity shall rise suddenly; and who non commiscearis 24:22 quoniam repente consurget perdi-

knoweth the ruin of them both? 24:23 These things also belong to the wise. It is not good tio eorum et ruinam utriusque quis nouit 24:23 haec quoto have respect of persons in judgment.

que sapientibus cognoscere personam in iudicio non est bo-

24:24 He that saith unto the wicked, Thou are righteous; num 24:24 qui dicit impio iustus es maledicent ei populi et

him shall the people curse, nations shall abhor him:

24:25 But to them that rebuke him shall be delight, and a detestabuntur eum tribus 24:25 qui arguunt laudabuntur et

good blessing shall come upon them.

super ipsos ueniet benedictio 24:26 labia deosculabitur qui 24:26 Every man shall kiss his lips that giveth a right an-

swer.

recta uerba respondet 24:27 praepara foris opus tuum et di- 24:27 Prepare thy work without, and make it fit for thyself

in the field; and afterwards build thine house.

ligenter exerce agrum tuum ut postea aedifices domum tuam 24:28 ne sis testis frustra contra proximum tuum nec lactes 24:28 Be not a witness against thy neighbour without

cause; and deceive not with thy lips.

quemquam labiis tuis 24:29 ne dicas quomodo fecit mihi 24:29 Say not, I will do so to him as he hath done to me: I

will render to the man according to his work.

sic faciam ei reddam unicuique secundum opus suum 24:30 24:30 I went by the field of the slothful, and by the vine-

yard of the man void of understanding;

per agrum hominis pigri transiui et per uineam uiri stulti 24:31 et ecce totum repleuerant urticae operuerant superfi- 24:31 And, lo, it was all grown over with thorns, and

nettles had covered the face thereof, and the stone wall thereof was broken down.

ciem eius spinae et maceria lapidum destructa erat 24:32

24:32 Then I saw, and considered it well: I looked upon it, and received instruction.

quod cum uidissem posui in corde meo et exemplo didici

disciplinam 24:33 parum inquam dormies modicum dormi- 24:33 Yet a little sleep, a little slumber, a little folding of

the hands to sleep:

tabis pauxillum manus conseres ut quiescas 24:34 et ueniet 24:34 So shall thy poverty come as one that travelleth; and

thy want as an armed man.

quasi cursor egestas tua et mendicitas quasi uir armatus

25:1 haec quoque parabolae Salomonis quas transtulerunt 25:1 These are also proverbs of Solomon, which the men

of Hezekiah king of Judah copied out.

uiri Ezechiae regis Iuda 25:2 gloria Dei celare uerbum et 25:2 It is the glory of God to conceal a thing: but the hon-

our of kings is to search out a matter.

gloria regum inuestigare sermonem 25:3 caelum sursum et 25:3 The heaven for height, and the earth for depth, and

the heart of kings is unsearchable.

terra deorsum et cor regum inscrutabile 25:4 aufer robi- 25:4 Take away the dross from the silver, and there shall

come forth a vessel for the finer.

ginem de argento et egredietur uas purissimum 25:5 au- 25:5 Take away the wicked from before the king, and his

throne shall be established in righteousness.

fer impietatem de uultu regis et firmabitur iustitia thronus

eius 25:6 ne gloriosus appareas coram rege et in loco ma- 25:6 Put not forth thyself in the presence of the king, and

stand not in the place of great men:

gnorum ne steteris 25:7 melius est enim ut dicatur tibi as- 25:7 For better it is that it be said unto thee, Come up

hither; than that thou shouldest be put lower in the presence of the prince whom thine eyes have seen.

cende huc quam ut humilieris coram principe 25:8 quae ui-

25:8 Go not forth hastily to strive, lest thou know not what to do in the end thereof, when thy neighbour hath put thee to shame.

derunt oculi tui ne proferas in iurgio cito ne postea emendare non possis cum dehonestaueris amicum tuum 25:9 causam 25:9 Debate thy cause with thy neighbour himself; and

discover not a secret to another:

tuam tracta cum amico tuo et secretum extraneo non reueles 25:10 ne forte insultet tibi cum audierit et exprobrare 25:10 Lest he that heareth it put thee to shame, and thine

infamy turn not away.

non cesset 25:11 mala aurea in lectis argenteis qui loquitur 25:11 A word fitly spoken is like apples of gold in pictures

of silver.

uerbum in tempore suo 25:12 inauris aurea et margaritum 25:12 As an earring of gold, and an ornament of fine gold,

so is a wise reprover upon an obedient ear.

fulgens qui arguit sapientem et aurem oboedientem 25:13 25:13 As the cold of snow in the time of harvest, so is a

faithful messenger to them that send him: for he refresheth the soul of his masters.

sicut frigus niuis in die messis ita legatus fidelis ei qui misit eum animam illius requiescere facit 25:14 nubes et uen- 25:14 Whoso boasteth himself of a false gift is like clouds

and wind without rain.

tus et pluuiae non sequentes uir gloriosus et promissa non conplens 25:15 patientia lenietur princeps et lingua mollis 25:15 By long forbearing is a prince persuaded, and a soft

tongue breaketh the bone.

25:16 Hast thou found honey? eat so much as is sufficient confringet duritiam 25:16 mel inuenisti comede quod suf-

for thee, lest thou be filled therewith, and vomit it.

25:17 Withdraw thy foot from thy neighbour’s house; lest ficit tibi ne forte saturatus euomas illud 25:17 subtrahe pe-

he be weary of thee, and so hate thee.

dem tuum de domo proximi tui nequando satiatus oderit te

25:18 A man that beareth false witness against his neigh- 25:18 iaculum et gladius et sagitta acuta homo qui loquitur

bour is a maul, and a sword, and a sharp arrow.

25:19 Confidence in an unfaithful man in time of trouble contra proximum suum testimonium falsum 25:19 dens pu-

is like a broken tooth, and a foot out of joint.

tridus et pes lapsus qui sperat super infideli in die angus-

25:20 As he that taketh away a garment in cold weather, tiae 25:20 et amittit pallium in die frigoris acetum in nitro et

and as vinegar upon nitre, so is he that singeth songs to an heavy heart.

qui cantat carmina cordi pessimo 25:21 si esurierit inimicus

25:21 If thine enemy be hungry, give him bread to eat; and if he be thirsty, give him water to drink:

25:22 For thou shalt heap coals of fire upon his head, and tuus ciba illum et si sitierit da ei aquam bibere 25:22 pru-

the LORD shall reward thee.

nam enim congregabis super caput eius et Dominus reddet

25:23 The north wind driveth away rain: so doth an angry tibi 25:23 uentus aquilo dissipat pluuias et facies tristis lin-

countenance a backbiting tongue.

25:24 It is better to dwell in the corner of the housetop, guam detrahentem 25:24 melius est sedere in angulo doma-

than with a brawling woman and in a wide house.

25:25 As cold waters to a thirsty soul, so is good news tis quam cum muliere litigiosa et in domo communi 25:25

from a far country.

aqua frigida animae sitienti et nuntius bonus de terra longin-

25:26 A righteous man falling down before the wicked is qua 25:26 fons turbatus pede et uena corrupta iustus cadens

as a troubled fountain, and a corrupt spring.

25:27 It is not good to eat much honey: so for men to coram impio 25:27 sicut qui mel multum comedit non est

search their own glory is not glory.

ei bonum sic qui scrutator est maiestatis opprimitur gloria

25:28 He that hath no rule over his own spirit is like a city 25:28 sicut urbs patens et absque murorum ambitu ita uir qui

that is broken down, and without walls.

non potest in loquendo cohibere spiritum suum

26:1 As snow in summer, and as rain in harvest, so honour 26:1 quomodo nix aestate et pluuia in messe sic indecens

is not seemly for a fool.

26:2 As the bird by wandering, as the swallow by flying, est stulto gloria 26:2 sicut auis ad alia transuolans et passer

so the curse causeless shall not come.

quolibet uadens sic maledictum frustra prolatum in quem-

26:3 A whip for the horse, a bridle for the ass, and a rod piam superueniet 26:3 flagellum equo et camus asino et for the fool’s back.

26:4 Answer not a fool according to his folly, lest thou uirga dorso inprudentium 26:4 ne respondeas stulto iuxta

also be like unto him.

26:5 Answer a fool according to his folly, lest he be wise stultitiam suam ne efficiaris ei similis 26:5 responde stulto

in his own conceit.

26:6 He that sendeth a message by the hand of a fool cut- iuxta stultitiam suam ne sibi sapiens esse uideatur 26:6

teth off the feet, and drinketh damage.

claudus pedibus et iniquitatem bibens qui mittit uerba per

26:7 The legs of the lame are not equal: so is a parable in nuntium stultum 26:7 quomodo pulchras frustra habet clau-

the mouth of fools.

26:8 As he that bindeth a stone in a sling, so is he that dus tibias sic indecens est in ore stultorum parabola 26:8

giveth honour to a fool.

sicut qui mittit lapidem in aceruum Mercurii ita qui tri-

26:9 As a thorn goeth up into the hand of a drunkard, so is buit insipienti honorem 26:9 quomodo si spina nascatur in

a parable in the mouths of fools. 26:10 The great God that formed all things both rewardeth manu temulenti sic parabola in ore stultorum 26:10 iudithe fool, and rewardeth transgressors.

cium determinat causas et qui inponit stulto silentium iras mitigat 26:11 sicut canis qui reuertitur ad uomitum suum 26:11 As a dog returneth to his vomit, so a fool returneth

to his folly.

sic inprudens qui iterat stultitiam suam 26:12 uidisti homi- 26:12 Seest thou a man wise in his own conceit? there is

more hope of a fool than of him.

nem sapientem sibi uideri magis illo spem habebit stultus 26:13 dicit piger leaena in uia leo in itineribus 26:14 sicut 26:13 The slothful man saith, There is a lion in the way; a

lion is in the streets.

26:14 As the door turneth upon his hinges, so doth the slothful upon his bed.

ostium uertitur in cardine suo ita piger in lectulo suo 26:15

26:15 The slothful hideth his hand in his bosom; it grieveth

abscondit piger manus sub ascellas suas et laborat si ad os ^{him to bring it again to his mouth.} suum eas conuerterit 26:16 sapientior sibi piger uidetur se- 26:16 The sluggard is wiser in his own conceit than seven

men that can render a reason.

ptem uiris loquentibus sententias 26:17 sicut qui adprehen- 26:17 He that passeth by, and meddleth with strife belong-

ing not to him, is like one that taketh a dog by the ears.

dit auribus canem sic qui transit et inpatiens commiscetur rixae alterius 26:18 sicut noxius est qui mittit lanceas et sa- 26:18 As a mad man who casteth firebrands, arrows, and

death,

gittas et mortem 26:19 sic uir qui fraudulenter nocet amico 26:19 So is the man that deceiveth his neighbour, and

saith, Am not I in sport?

suo et cum fuerit deprehensus dicit ludens feci 26:20 cum 26:20 Where no wood is, there the fire goeth out: so where

there is no talebearer, the strife ceaseth.

defecerint ligna extinguetur ignis et susurrone subtracto iurgia conquiescunt 26:21 sicut carbones ad prunam et ligna 26:21 As coals are to burning coals, and wood to fire; so

is a contentious man to kindle strife.

ad ignem sic homo iracundus suscitat rixas 26:22 uerba su- 26:22 The words of a talebearer are as wounds, and they

go down into the innermost parts of the belly.

surronis quasi simplicia et ipsa perueniunt ad intima uentris 26:23 quomodo si argento sordido ornare uelis uas fictile sic 26:23 Burning lips and a wicked heart are like a potsherd

covered with silver dross.

labia tumentia cum pessimo corde sociata 26:24 labiis suis 26:24 He that hateth dissembleth with his lips, and layeth

up deceit within him;

intellegitur inimicus cum in corde tractauerit dolos 26:25 26:25 When he speaketh fair, believe him not: for there

are seven abominations in his heart.

quando submiserit uocem suam ne credideris ei quoniam septem nequitiae sunt in corde illius 26:26 qui operit odium 26:26 Whose hatred is covered by deceit, his wickedness

shall be shewed before the whole congregation.

fraudulenter reuelabitur malitia eius in concilio 26:27 qui 26:27 Whoso diggeth a pit shall fall therein: and he that

rolleth a stone, it will return upon him.

fodit foueam incidet in eam et qui uoluit lapidem reuertetur ad eum 26:28 lingua fallax non amat ueritatem et os lubri- 26:28 A lying tongue hateth those that are afflicted by it;

and a flattering mouth worketh ruin.

cum operatur ruinas

27:1 ne glorieris in crastinum ignorans quid superuentura 27:1 Boast not thyself of to morrow; for thou knowest not

what a day may bring forth.

pariat dies 27:2 laudet te alienus et non os tuum extraneus 27:2 Let another man praise thee, and not thine own

mouth; a stranger, and not thine own lips.

et non labia tua 27:3 graue est saxum et onerosa harena sed 27:3 A stone is heavy, and the sand weighty; but a fool’s

wrath is heavier than them both.

ira stulti utroque grauior 27:4 ira non habet misericordiam 27:4 Wrath is cruel, and anger is outrageous; but who is

able to stand before envy?

nec erumpens furor et impetum concitati ferre quis poterit 27:5 melior est manifesta correptio quam amor absconditus 27:5 Open rebuke is better than secret love. 27:6 meliora sunt uulnera diligentis quam fraudulenta odien- 27:6 Faithful are the wounds of a friend; but the kisses of

an enemy are deceitful.

tis oscula 27:7 anima saturata calcabit fauum anima esu- 27:7 The full soul loatheth an honeycomb; but to the hun-

gry soul every bitter thing is sweet.

27:8 As a bird that wandereth from her nest, so is a man riens et amarum pro dulce sumet 27:8 sicut auis transmig-

that wandereth from his place. 27:9 Ointment and perfume rejoice the heart: so doth the rans de nido suo sic uir qui relinquit locum suum 27:9 unsweetness of a man’s friend by hearty counsel.

guento et uariis odoribus delectatur cor et bonis amici consi-

27:10 Thine own friend, and thy father’s friend, forsake liis anima dulcoratur 27:10 amicum tuum et amicum patris not; neither go into thy brother’s house in the day of thy calamity: for better is a neighbour that is near than a

brother far off. tui ne dimiseris et domum fratris tui ne ingrediaris in die adflictionis tuae melior est uicinus iuxta quam frater pro-

27:11 My son, be wise, and make my heart glad, that I cul 27:11 stude sapientiae fili mi et laetifica cor meum ut

may answer him that reproacheth me.

27:12 A prudent man foreseeth the evil, and hideth him- possim exprobranti respondere sermonem 27:12 astutus ui-

self; but the simple pass on, and are punished.

dens malum absconditus est paruuli transeuntes sustinuere

27:13 Take his garment that is surety for a stranger, and dispendia 27:13 tolle uestimentum eius qui spopondit pro

take a pledge of him for a strange woman.

27:14 He that blesseth his friend with a loud voice, rising extraneo et pro alienis auferto pignus 27:14 qui benedicit

early in the morning, it shall be counted a curse to him.

proximo suo uoce grandi de nocte consurgens maledicenti

27:15 A continual dropping in a very rainy day and a con- similis erit 27:15 tecta perstillantia in die frigoris et litigiosa

tentious woman are alike.

27:16 Whosoever hideth her hideth the wind, and the oint- mulier conparantur 27:16 qui retinet eam quasi qui uentum

ment of his right hand, which bewrayeth itself. 27:17 Iron sharpeneth iron; so a man sharpeneth the coun- teneat et oleum dexterae suae uocabit 27:17 ferrum ferro tenance of his friend. 27:18 Whoso keepeth the fig tree shall eat the fruit thereof: acuitur et homo exacuit faciem amici sui 27:18 qui seruat so he that waiteth on his master shall be honoured.

ficum comedet fructus eius et qui custos est domini sui glo-

27:19 As in water face answereth to face, so the heart of rificabitur 27:19 quomodo in aquis resplendent uultus pro-

man to man.

spicientium sic corda hominum manifesta sunt prudentibus

27:20 Hell and destruction are never full; so the eyes of 27:20 infernus et perditio non replentur similiter et oculi ho-

man are never satisfied. 27:21 As the fining pot for silver, and the furnace for gold; minum insatiabiles 27:21 quomodo probatur in conflatorio so is a man to his praise.

argentum et in fornace aurum sic probatur homo ore laudan-

27:22 Though thou shouldest bray a fool in a mortar tis 27:22 si contuderis stultum in pila quasi tisanas feriente

among wheat with a pestle, yet will not his foolishness depart from him.

desuper pilo non auferetur ab eo stultitia eius 27:23 dili-

27:23 Be thou diligent to know the state of thy flocks, and look well to thy herds.

genter agnosce uultum pecoris tui tuosque greges considera

27:24 For riches are not for ever: and doth the crown en- 27:24 non enim habebis iugiter potestatem sed corona tribue-

dure to every generation?

27:25 The hay appeareth, and the tender grass sheweth tur in generatione generationum 27:25 aperta sunt prata et

itself, and herbs of the mountains are gathered.

apparuerunt herbae uirentes et collecta sunt faena de mon-

27:26 The lambs are for thy clothing, and the goats are the tibus 27:26 agni ad uestimentum tuum et hedi agri pretium price of the field.

27:27 And thou shalt have goats’ milk enough for thy 27:27 sufficiat tibi lac caprarum in cibos tuos in necessaria

food, for the food of thy household, and for the maintenance for thy maidens.

domus tuae et ad uictum ancillis tuis

28:1 The wicked flee when no man pursueth: but the righ- 28:1 fugit impius nemine persequente iustus autem quasi

teous are bold as a lion.

28:2 For the transgression of a land many are the princes leo confidens absque terrore erit 28:2 propter peccata ter-

thereof: but by a man of understanding and knowledge the state thereof shall be prolonged.

rae multi principes eius et propter hominis sapientiam et horum scientiam quae dicuntur uita ducis longior erit 28:3 uir 28:3 A poor man that oppresseth the poor is like a sweep-

ing rain which leaveth no food.

pauper calumnians pauperes similis imbri uehementi in quo paratur fames 28:4 qui derelinquunt legem laudant impium 28:4 They that forsake the law praise the wicked: but such

as keep the law contend with them.

qui custodiunt succenduntur contra eum 28:5 uiri mali non 28:5 Evil men understand not judgment: but they that seek

the LORD understand all things.

cogitant iudicium qui autem requirunt Dominum animaduer-

tunt omnia 28:6 melior est pauper ambulans in simplicitate 28:6 Better is the poor that walketh in his uprightness, than

he that is perverse in his ways, though he be rich.

sua quam diues prauis itineribus 28:7 qui custodit legem 28:7 Whoso keepeth the law is a wise son: but he that is a

companion of riotous men shameth his father.

filius sapiens est qui pascit comesatores confundit patrem suum 28:8 qui coaceruat diuitias usuris et fenore liberali in 28:8 He that by usury and unjust gain increaseth his sub-

stance, he shall gather it for him that will pity the poor.

pauperes congregat eas 28:9 qui declinat aurem suam ne 28:9 He that turneth away his ear from hearing the law,

even his prayer shall be abomination.

audiat legem oratio eius erit execrabilis 28:10 qui decipit 28:10 Whoso causeth the righteous to go astray in an evil

way, he shall fall himself into his own pit: but the upright shall have good things in possession.

iustos in uia mala in interitu suo corruet et simplices possidebunt bona 28:11 sapiens sibi uidetur uir diues pauper 28:11 The rich man is wise in his own conceit; but the

poor that hath understanding searcheth him out.

autem prudens scrutabitur eum 28:12 in exultatione iusto- 28:12 When righteous men do rejoice, there is great glory:

but when the wicked rise, a man is hidden.

rum multa gloria regnantibus impiis ruinae hominum 28:13 28:13 He that covereth his sins shall not prosper: but

whoso confesseth and forsaketh them shall have mercy.

qui abscondit scelera sua non dirigetur qui confessus fuerit et reliquerit ea misericordiam consequetur 28:14 beatus 28:14 Happy is the man that feareth alway: but he that

hardeneth his heart shall fall into mischief.

homo qui semper est pauidus qui uero mentis est durae corruet in malum 28:15 leo rugiens et ursus esuriens princeps 28:15 As a roaring lion, and a ranging bear; so is a wicked

ruler over the poor people.

impius super populum pauperem 28:16 dux indigens pru- 28:16 The prince that wanteth understanding is also a great oppressor: but he that hateth covetousness shall prolong his days.

dentia multos opprimet per calumniam qui autem odit auaritiam longi fient dies eius 28:17 hominem qui calumnia- 28:17 A man that doeth violence to the blood of any per-

son shall flee to the pit; let no man stay him.

tur animae sanguinem si usque ad lacum fugerit nemo sustentet 28:18 qui ambulat simpliciter saluus erit qui peruer- 28:18 Whoso walketh uprightly shall be saved: but he that

is perverse in his ways shall fall at once.

sis ingreditur uiis concidet semel 28:19 qui operatur terram 28:19 He that tilleth his land shall have plenty of bread:

but he that followeth after vain persons shall have poverty enough.

suam saturabitur panibus qui sectatur otium replebitur egestate 28:20 uir fidelis multum laudabitur qui autem festinat 28:20 A faithful man shall abound with blessings: but he

that maketh haste to be rich shall not be innocent.

ditari non erit innocens 28:21 qui cognoscit in iudicio fa- 28:21 To have respect of persons is not good: for for a

piece of bread that man will transgress.

ciem non facit bene iste et pro buccella panis deserit ueritatem 28:22 uir qui festinat ditari et aliis inuidet ignorat quod 28:22 He that hasteth to be rich hath an evil eye, and con-

sidereth not that poverty shall come upon him.

egestas superueniat ei 28:23 qui corripit hominem gratiam 28:23 He that rebuketh a man afterwards shall find more

favour than he that flattereth with the tongue.

postea inueniet apud eum magis quam ille qui per linguae 28:24 Whoso robbeth his father or his mother, and saith, blandimenta decipit 28:24 qui subtrahit aliquid a patre suo

It is no transgression; the same is the companion of a destroyer.

et matre et dicit hoc non est peccatum particeps homicidae

28:25 He that is of a proud heart stirreth up strife: but he est 28:25 qui se iactat et dilatat iurgia concitat qui sperat

that putteth his trust in the LORD shall be made fat.

28:26 He that trusteth in his own heart is a fool: but whoso in Domino saginabitur 28:26 qui confidit in corde suo stul-

walketh wisely, he shall be delivered.

28:27 He that giveth unto the poor shall not lack: but he tus est qui autem graditur sapienter iste saluabitur 28:27 qui

that hideth his eyes shall have many a curse.

dat pauperi non indigebit qui despicit deprecantem sustinebit

28:28 When the wicked rise, men hide themselves: but penuriam 28:28 cum surrexerint impii abscondentur homi-

when they perish, the righteous increase.

nes cum illi perierint multiplicabuntur iusti

29:1 He, that being often reproved hardeneth his neck, 29:1 uiro qui corripientem dura ceruice contemnit repenti-

shall suddenly be destroyed, and that without remedy.

29:2 When the righteous are in authority, the people re- nus superueniet interitus et eum sanitas non sequitur 29:2 in

joice: but when the wicked beareth rule, the people mourn.

multiplicatione iustorum laetabitur uulgus cum impii sum-

29:3 Whoso loveth wisdom rejoiceth his father: but he that pserint principatum gemet populus 29:3 uir qui amat sakeepeth company with harlots spendeth his substance.

pientiam laetificat patrem suum qui autem nutrit scorta per-

29:4 The king by judgment establisheth the land: but he det substantiam 29:4 rex iustus erigit terram uir auarus de-

that receiveth gifts overthroweth it.

29:5 A man that flattereth his neighbour spreadeth a net struet eam 29:5 homo qui blandis fictisque sermonibus lo-

for his feet.

29:6 In the transgression of an evil man there is a snare: quitur amico suo rete expandit gressibus eius 29:6 peccan-

but the righteous doth sing and rejoice.

tem uirum iniquum inuoluet laqueus et iustus laudabit at-

29:7 The righteous considereth the cause of the poor: but que gaudebit 29:7 nouit iustus causam pauperum impius the wicked regardeth not to know it.

29:8 Scornful men bring a city into a snare: but wise men ignorat scientiam 29:8 homines pestilentes dissipant ciui-

turn away wrath.

29:9 If a wise man contendeth with a foolish man, whether tatem sapientes auertunt furorem 29:9 uir sapiens si cum

he rage or laugh, there is no rest.

stulto contenderit siue irascatur siue rideat non inueniet re-

29:10 The bloodthirsty hate the upright: but the just seek quiem 29:10 uiri sanguinum oderunt simplicem iusti quaehis soul.

29:11 A fool uttereth all his mind: but a wise man keepeth runt animam eius 29:11 totum spiritum suum profert stultus

it in till afterwards. 29:12 If a ruler hearken to lies, all his servants are wicked. sapiens differt et reseruat in posterum 29:12 princeps qui

libenter audit uerba mendacii omnes ministros habebit im-

29:13 The poor and the deceitful man meet together: the pios 29:13 pauper et creditor obuiam fuerunt sibi utriusque

LORD lighteneth both their eyes. 29:14 The king that faithfully judgeth the poor, his throne inluminator est Dominus 29:14 rex qui iudicat in ueritate shall be established for ever.

29:15 The rod and reproof give wisdom: but a child left to pauperes thronus eius in aeternum firmabitur 29:15 uirga

himself bringeth his mother to shame.

atque correptio tribuet sapientiam puer autem qui dimitti-

29:16 When the wicked are multiplied, transgression in- tur uoluntati suae confundet matrem suam 29:16 in multicreaseth: but the righteous shall see their fall.

plicatione impiorum multiplicabuntur scelera et iusti ruinas

29:17 Correct thy son, and he shall give thee rest; yea, he eorum uidebunt 29:17 erudi filium tuum et refrigerabit te shall give delight unto thy soul.

29:18 Where there is no vision, the people perish: but he et dabit delicias animae tuae 29:18 cum prophetia defece-

that keepeth the law, happy is he.

rit dissipabitur populus qui custodit legem beatus est 29:19 29:19 A servant will not be corrected by words: for though

he understand he will not answer.

seruus uerbis non potest erudiri quia quod dicis intellegit et respondere contemnit 29:20 uidisti hominem uelocem ad 29:20 Seest thou a man that is hasty in his words? there is

more hope of a fool than of him.

loquendum stulti magis speranda est quam illius correptio 29:21 qui delicate a pueritia nutrit seruum suum postea il- 29:21 He that delicately bringeth up his servant from a

child shall have him become his son at the length.

lum sentiet contumacem 29:22 uir iracundus prouocat rixas 29:22 An angry man stirreth up strife, and a furious man

aboundeth in transgression.

et qui ad indignandum facilis est erit ad peccata procliuior 29:23 superbum sequitur humilitas et humilem spiritu susci- 29:23 A man’s pride shall bring him low: but honour shall

uphold the humble in spirit.

piet gloria 29:24 qui cum fure partitur odit animam suam 29:24 Whoso is partner with a thief hateth his own soul:

he heareth cursing, and bewrayeth it not.

adiurantem audit et non indicat 29:25 qui timet hominem 29:25 The fear of man bringeth a snare: but whoso putteth

his trust in the LORD shall be safe.

cito corruet qui sperat in Domino subleuabitur 29:26 multi 29:26 Many seek the ruler’s favour; but every man’s judg-

ment cometh from the LORD.

requirunt faciem principis et a Domino iudicium egreditur singulorum 29:27 abominantur iusti uirum impium et abo- 29:27 An unjust man is an abomination to the just: and he

that is upright in the way is abomination to the wicked.

minantur impii eos qui in recta sunt uia

30:1 uerba Congregantis filii Vomentis uisio quam locutus 30:1 The words of Agur the son of Jakeh, even the

prophecy: the man spake unto Ithiel, even unto Ithiel and Ucal,

est uir cum quo est Deus et qui Deo secum morante confortatus ait 30:2 stultissimus sum uirorum et sapientia ho- 30:2 Surely I am more brutish than any man, and have not

the understanding of a man.

minum non est mecum 30:3 non didici sapientiam et non 30:3 I neither learned wisdom, nor have the knowledge of

the holy.

noui sanctorum scientiam 30:4 quis ascendit in caelum at- 30:4 Who hath ascended up into heaven, or descended?

who hath gathered the wind in his fists? who hath bound the waters in a garment? who hath established all the ends of the earth? what is his name, and what is his son’s name, if thou canst tell?

que descendit quis continuit spiritum manibus suis quis conligauit aquas quasi in uestimento quis suscitauit omnes terminos terrae quod nomen eius et quod nomen filii eius si nosti 30:5 omnis sermo Dei ignitus clypeus est speranti- 30:5 Every word of God is pure: he is a shield unto them

that put their trust in him.

bus in se 30:6 ne addas quicquam uerbis illius et arguaris 30:6 Add thou not unto his words, lest he reprove thee,

and thou be found a liar.

inueniarisque mendax 30:7 duo rogaui te ne deneges mihi 30:7 Two things have I required of thee; deny me them not

before I die:

antequam moriar 30:8 uanitatem et uerba mendacia longe 30:8 Remove far from me vanity and lies: give me neither

poverty nor riches; feed me with food convenient for me:

fac a me mendicitatem et diuitias ne dederis mihi tribue tantum uictui meo necessaria 30:9 ne forte saturatus inliciar ad 30:9 Lest I be full, and deny thee, and say, Who is the

LORD? or lest I be poor, and steal, and take the name of my God in vain.

negandum et dicam quis est Dominus et egestate conpulsus

furer et peierem nomen Dei mei 30:10 ne accuses seruum 30:10 Accuse not a servant unto his master, lest he curse

thee, and thou be found guilty.

ad dominum suum ne forte maledicat tibi et corruas 30:11 30:11 There is a generation that curseth their father, and

doth not bless their mother.

generatio quae patri suo maledicit et quae non benedicit ma-

tri suae 30:12 generatio quae sibi munda uidetur et tamen 30:12 There is a generation that are pure in their own eyes, and yet is not washed from their filthiness.

30:13 There is a generation, O how lofty are their eyes! non est lota a sordibus suis 30:13 generatio cuius excelsi

and their eyelids are lifted up.

30:14 There is a generation, whose teeth are as swords, sunt oculi et palpebrae eius in alta subrectae 30:14 genera-

and their jaw teeth as knives, to devour the poor from off the earth, and the needy from among men.

tio quae pro dentibus gladios habet et commandit molaribus suis ut comedat inopes de terra et pauperes ex hominibus

30:15 The horseleach hath two daughters, crying, Give, 30:15 sanguisugae duae sunt filiae dicentes adfer adfer tria

give. There are three things that are never satisfied, yea, four things say not, It is enough:

sunt insaturabilia et quartum quod numquam dicit sufficit

30:16 The grave; and the barren womb; the earth that is not 30:16 infernus et os uuluae et terra quae non satiatur aqua ig-

filled with water; and the fire that saith not, It is enough. 30:17 The eye that mocketh at his father, and despiseth to nis uero numquam dicit sufficit 30:17 oculum qui subsanobey his mother, the ravens of the valley shall pick it out, and the young eagles shall eat it.

nat patrem et qui despicit partum matris suae effodiant corui

30:18 There be three things which are too wonderful for de torrentibus et comedant illum filii aquilae 30:18 tria sunt

me, yea, four which I know not:

30:19 The way of an eagle in the air; the way of a serpent difficilia mihi et quartum penitus ignoro 30:19 uiam aqui-

upon a rock; the way of a ship in the midst of the sea; and the way of a man with a maid.

lae in caelo uiam colubri super petram uiam nauis in medio

30:20 Such is the way of an adulterous woman; she eateth, mari et uiam uiri in adulescentula 30:20 talis est uia mulie-

and wipeth her mouth, and saith, I have done no wickedness.

ris adulterae quae comedit et tergens os suum dicit non sum

30:21 For three things the earth is disquieted, and for four operata malum 30:21 per tria mouetur terra et quartum non which it cannot bear:

30:22 For a servant when he reigneth; and a fool when he potest sustinere 30:22 per seruum cum regnauerit per stul-

is filled with meat;

30:23 For an odious woman when she is married; and an tum cum saturatus fuerit cibo 30:23 per odiosam mulierem

handmaid that is heir to her mistress.

cum in matrimonio fuerit adsumpta et per ancillam cum he-

30:24 There be four things which are little upon the earth, res fuerit dominae suae 30:24 quattuor sunt minima terrae but they are exceeding wise:

30:25 The ants are a people not strong, yet they prepare et ipsa sunt sapientiora sapientibus 30:25 formicae popu-

their meat in the summer; 30:26 The conies are but a feeble folk, yet make they their lus infirmus quae praeparant in messe cibum sibi 30:26 lehouses in the rocks;

pusculus plebs inualida quae conlocat in petra cubile suum

30:27 The locusts have no king, yet go they forth all of 30:27 regem lucusta non habet et egreditur uniuersa per tur-

them by bands;

30:28 The spider taketh hold with her hands, and is in mas 30:28 stilio manibus nititur et moratur in aedibus regis

kings’ palaces.

30:29 There be three things which go well, yea, four are 30:29 tria sunt quae bene gradiuntur et quartum quod incedit

comely in going:

30:30 A lion which is strongest among beasts, and turneth feliciter 30:30 leo fortissimus bestiarum ad nullius pauebit

not away for any;

30:31 A greyhound; an he goat also; and a king, against occursum 30:31 gallus succinctus lumbos et aries nec est

whom there is no rising up. 30:32 If thou hast done foolishly in lifting up thyself, or if rex qui resistat ei 30:32 et qui stultus apparuit postquam thou hast thought evil, lay thine hand upon thy mouth.

elatus est in sublime si enim intellexisset ori inposuisset ma-

30:33 Surely the churning of milk bringeth forth butter, num 30:33 qui autem fortiter premit ubera ad eliciendum and the wringing of the nose bringeth forth blood: so the forcing of wrath bringeth forth strife.

lac exprimit butyrum et qui uehementer emungitur elicit sanguinem et qui prouocat iras producit discordias

31:1 The words of king Lemuel, the prophecy that his 31:1 uerba Lamuhel regis uisio qua erudiuit eum mater sua

mother taught him.

31:2 quid dilecte mi quid dilecte uteri mei quid dilecte uoto- 31:2 What, my son? and what, the son of my womb? and

what, the son of my vows?

rum meorum 31:3 ne dederis mulieribus substantiam tuam 31:3 Give not thy strength unto women, nor thy ways to

that which destroyeth kings.

et uias tuas ad delendos reges 31:4 noli regibus o Lamu- 31:4 It is not for kings, O Lemuel, it is not for kings to

drink wine; nor for princes strong drink:

hel noli regibus dare uinum quia nullum secretum est ubi regnat ebrietas 31:5 ne forte bibat et obliuiscatur iudicio- 31:5 Lest they drink, and forget the law, and pervert the

judgment of any of the afflicted.

rum et mutet causam filiorum pauperis 31:6 date siceram 31:6 Give strong drink unto him that is ready to perish,

and wine unto those that be of heavy hearts.

maerentibus et uinum his qui amaro sunt animo 31:7 bi- 31:7 Let him drink, and forget his poverty, and remember

his misery no more.

bant ut obliuiscantur egestatis suae et doloris non recordentur amplius 31:8 aperi os tuum muto et causis omnium fi- 31:8 Open thy mouth for the dumb in the cause of all such

as are appointed to destruction.

liorum qui pertranseunt 31:9 aperi os tuum decerne quod 31:9 Open thy mouth, judge righteously, and plead the

cause of the poor and needy.

iustum est et iudica inopem et pauperem 31:10 aleph mu- 31:10 Who can find a virtuous woman? for her price is far

above rubies.

lierem fortem quis inueniet procul et de ultimis finibus pretium eius 31:11 beth confidit in ea cor uiri sui et spoliis non 31:11 The heart of her husband doth safely trust in her, so

that he shall have no need of spoil.

indigebit 31:12 gimel reddet ei bonum et non malum omni- 31:12 She will do him good and not evil all the days of her

life.

bus diebus uitae suae 31:13 deleth quaesiuit lanam et linum 31:13 She seeketh wool, and flax, and worketh willingly

with her hands.

et operata est consilio manuum suarum 31:14 he facta est 31:14 She is like the merchants’ ships; she bringeth her

food from afar.

quasi nauis institoris de longe portat panem suum 31:15 uau 31:15 She riseth also while it is yet night, and giveth meat

to her household, and a portion to her maidens.

et de nocte surrexit deditque praedam domesticis suis et cibaria ancillis suis 31:16 zai considerauit agrum et emit eum 31:16 She considereth a field, and buyeth it: with the fruit

of her hands she planteth a vineyard.

de fructu manuum suarum plantauit uineam 31:17 heth ac- 31:17 She girdeth her loins with strength, and strength-

eneth her arms.

cinxit fortitudine lumbos suos et roborauit brachium suum 31:18 teth gustauit quia bona est negotiatio eius non extin- 31:18 She perceiveth that her merchandise is good: her

candle goeth not out by night.

guetur in nocte lucerna illius 31:19 ioth manum suam mi- 31:19 She layeth her hands to the spindle, and her hands

hold the distaff.

sit ad fortia et digiti eius adprehenderunt fusum 31:20 caph 31:20 She stretcheth out her hand to the poor; yea, she

reacheth forth her hands to the needy.

manum suam aperuit inopi et palmas suas extendit ad pauperem 31:21 lameth non timebit domui suae a frigoribus niuis 31:21 She is not afraid of the snow for her household: for

all her household are clothed with scarlet.

omnes enim domestici eius uestiti duplicibus 31:22 mem 31:22 She maketh herself coverings of tapestry; her cloth-

ing is silk and purple.

stragulam uestem fecit sibi byssus et purpura indumentum eius 31:23 nun nobilis in portis uir eius quando sederit cum 31:23 Her husband is known in the gates, when he sitteth

among the elders of the land.

senatoribus terrae 31:24 samech sindonem fecit et uendi- 31:24 She maketh fine linen, and selleth it; and delivereth

girdles unto the merchant.

dit et cingulum tradidit Chananeo 31:25 ain fortitudo et de- 31:25 Strength and honour are her clothing; and she shall

rejoice in time to come.

cor indumentum eius et ridebit in die nouissimo 31:26 phe 31:26 She openeth her mouth with wisdom; and in her

tongue is the law of kindness.

os suum aperuit sapientiae et lex clementiae in lingua eius 31:27 She looketh well to the ways of her household, and 31:27 sade considerat semitas domus suae et panem otiosa

eateth not the bread of idleness. 31:28 Her children arise up, and call her blessed; her hus- non comedet 31:28 coph surrexerunt filii eius et beatissiband also, and he praiseth her.

31:29 Many daughters have done virtuously, but thou ex- mam praedicauerunt uir eius et laudauit eam 31:29 res mul-

cellest them all.

tae filiae congregauerunt diuitias tu supergressa es uniuersas

31:30 Favour is deceitful, and beauty is vain: but a woman 31:30 sin fallax gratia et uana est pulchritudo mulier timens

that feareth the LORD, she shall be praised.

31:31 Give her of the fruit of her hands; and let her own Dominum ipsa laudabitur 31:31 thau date ei de fructu ma-

works praise her in the gates.

nuum suarum et laudent eam in portis opera eius

Liber Ecclesiastes

1:1 uerba Ecclesiastes filii Dauid regis Hierusalem 1:2 1:1 The words of the Preacher, the son of David, king in

Jerusalem.

1:2 Vanity of vanities, saith the Preacher, vanity of vanities; all is vanity.

uanitas uanitatum dixit Ecclesiastes uanitas uanitatum omnia uanitas 1:3 quid habet amplius homo de uniuerso labore 1:3 What profit hath a man of all his labour which he

taketh under the sun?

suo quod laborat sub sole 1:4 generatio praeterit et genera- 1:4 One generation passeth away, and another generation

cometh: but the earth abideth for ever.

tio aduenit terra uero in aeternum stat 1:5 oritur sol et occi- 1:5 The sun also ariseth, and the sun goeth down, and

hasteth to his place where he arose.

dit et ad locum suum reuertitur ibique renascens 1:6 gyrat 1:6 The wind goeth toward the south, and turneth about

unto the north; it whirleth about continually, and the wind returneth again according to his circuits.

per meridiem et flectitur ad aquilonem lustrans uniuersa circuitu pergit spiritus et in circulos suos regreditur 1:7 omnia 1:7 All the rivers run into the sea; yet the sea is not full;

unto the place from whence the rivers come, thither they return again.

flumina intrant mare et mare non redundat ad locum unde exeunt flumina reuertuntur ut iterum fluant 1:8 cunctae res 1:8 All things are full of labour; man cannot utter it: the

eye is not satisfied with seeing, nor the ear filled with hearing.

difficiles non potest eas homo explicare sermone non saturatur oculus uisu nec auris impletur auditu 1:9 quid est quod 1:9 The thing that hath been, it is that which shall be; and

that which is done is that which shall be done: and there is no new thing under the sun.

fuit ipsum quod futurum est quid est quod factum est ipsum

quod fiendum est 1:10 nihil sub sole nouum nec ualet quis- 1:10 Is there any thing whereof it may be said, See, this is new? it hath been already of old time, which was before us.

quam dicere ecce hoc recens est iam enim praecessit in saeculis quae fuerunt ante nos 1:11 non est priorum memoria 1:11 There is no remembrance of former things; neither

shall there be any remembrance of things that are to come with those that shall come after.

sed nec eorum quidem quae postea futura sunt erit recordatio apud eos qui futuri sunt in nouissimo 1:12 ego Ecclesiastes 1:12 I the Preacher was king over Israel in Jerusalem. fui rex Israhel in Hierusalem 1:13 et proposui in animo meo 1:13 And I gave my heart to seek and search out by wis-

dom concerning all things that are done under heaven: this sore travail hath God given to the sons of man to be exer-

quaerere et inuestigare sapienter de omnibus quae fiunt sub cised therewith. sole hanc occupationem pessimam dedit Deus filiis homi-

num ut occuparentur in ea 1:14 uidi quae fiunt cuncta sub 1:14 I have seen all the works that are done under the sun;

and, behold, all is vanity and vexation of spirit.

sole et ecce uniuersa uanitas et adflictio spiritus 1:15 per- 1:15 That which is crooked cannot be made straight: and

that which is wanting cannot be numbered.

uersi difficile corriguntur et stultorum infinitus est numerus 1:16 locutus sum in corde meo dicens ecce magnus effec- 1:16 I communed with mine own heart, saying, Lo, I am

come to great estate, and have gotten more wisdom than all they that have been before me in Jerusalem: yea, my

tus sum et praecessi sapientia omnes qui fuerunt ante me in heart had great experience of wisdom and knowledge. Hierusalem et mens mea contemplata est multa sapienter et didicit 1:17 dedique cor meum ut scirem prudentiam atque 1:17 And I gave my heart to know wisdom, and to know

madness and folly: I perceived that this also is vexation of spirit.

doctrinam erroresque et stultitiam et agnoui quod in his quoque esset labor et adflictio spiritus 1:18 eo quod in multa 1:18 For in much wisdom is much grief: and he that in-

creaseth knowledge increaseth sorrow.

sapientia multa sit indignatio et qui addit scientiam addat et laborem

2:1 dixi ego in corde meo uadam et affluam deliciis et fruar 2:1 I said in mine heart, Go to now, I will prove thee with

mirth, therefore enjoy pleasure: and, behold, this also is vanity.

Biblia Sacra Vulgata

2:2 I said of laughter, It is mad: and of mirth, What doeth bonis et uidi quod hoc quoque esset uanitas 2:2 risum reit?

2:3 I sought in mine heart to give myself unto wine, yet putaui errorem et gaudio dixi quid frustra deciperis 2:3 co-

acquainting mine heart with wisdom; and to lay hold on folly, till I might see what was that good for the sons of men, which they should do under the heaven all the days of their life.

gitaui in corde meo abstrahere a uino carnem meam ut animum meum transferrem ad sapientiam deuitaremque stultitiam donec uiderem quid esset utile filiis hominum quod

2:4 I made me great works; I builded me houses; I planted facto opus est sub sole numero dierum uitae suae 2:4 mame vineyards:

gnificaui opera mea aedificaui mihi domos plantaui uineas

2:5 I made me gardens and orchards, and I planted trees in 2:5 feci hortos et pomeria et conseui ea cuncti generis arbo-

them of all kind of fruits:

2:6 I made me pools of water, to water therewith the wood ribus 2:6 extruxi mihi piscinas aquarum ut inrigarem sil-

that bringeth forth trees:

2:7 I got me servants and maidens, and had servants born uam lignorum germinantium 2:7 possedi seruos et ancil-

in my house; also I had great possessions of great and small cattle above all that were in Jerusalem before me:

las multamque familiam habui armenta quoque et magnos ouium greges ultra omnes qui fuerunt ante me in Hierusalem

2:8 I gathered me also silver and gold, and the peculiar 2:8 coaceruaui mihi argentum et aurum et substantias regum

treasure of kings and of the provinces: I gat me men singers and women singers, and the delights of the sons

of men, as musical instruments, and that of all sorts. ac prouinciarum feci mihi cantores et cantrices et delicias

filiorum hominum scyphos et urceos in ministerio ad uina

2:9 So I was great, and increased more than all that were fundenda 2:9 et supergressus sum opibus omnes qui fuebefore me in Jerusalem: also my wisdom remained with me.

runt ante me in Hierusalem sapientia quoque perseuerauit

2:10 And whatsoever mine eyes desired I kept not from mecum 2:10 et omnia quae desiderauerunt oculi mei non

them, I withheld not my heart from any joy; for my heart rejoiced in all my labour: and this was my portion of all

my labour. negaui eis nec prohibui cor quin omni uoluptate frueretur et

oblectaret se in his quae paraueram et hanc ratus sum partem

2:11 Then I looked on all the works that my hands had meam si uterer labore meo 2:11 cumque me conuertissem

wrought, and on the labour that I had laboured to do: and, behold, all was vanity and vexation of spirit, and there was

no profit under the sun. ad uniuersa opera quae fecerant manus meae et ad labores in

quibus frustra sudaueram uidi in omnibus uanitatem et ad-

2:12 And I turned myself to behold wisdom, and madness, flictionem animi et nihil permanere sub sole 2:12 transiui

and folly: for what can the man do that cometh after the king? even that which hath been already done.

ad contemplandam sapientiam erroresque et stultitiam quid

2:13 Then I saw that wisdom excelleth folly, as far as light est inquam homo ut sequi possit regem factorem suum 2:13 excelleth darkness.

et uidi quia tantum praecederet sapientia stultitiam quantum

2:14 The wise man’s eyes are in his head; but the fool differt lux tenebris 2:14 sapientis oculi in capite eius stul-

walketh in darkness: and I myself perceived also that one event happeneth to them all.

tus in tenebris ambulat et didici quod unus utriusque esset

2:15 Then said I in my heart, As it happeneth to the fool, interitus 2:15 et dixi in corde meo si unus et stulti et meus

so it happeneth even to me; and why was I then more wise? Then I said in my heart, that this also is vanity.

occasus erit quid mihi prodeet quod maiorem sapientiae dedi operam locutusque cum mente mea animaduerti quod hoc

2:16 For there is no remembrance of the wise more than quoque esset uanitas 2:16 non enim erit memoria sapientis

of the fool for ever; seeing that which now is in the days to come shall all be forgotten. And how dieth the wise man? as the fool.

Liber Ecclesiastes

similiter ut stulti in perpetuum et futura tempora obliuione cuncta pariter obruent moritur doctus similiter et indoctus 2:17 et idcirco taeduit me uitae meae uidentem mala esse uni- 2:17 Therefore I hated life; because the work that is

wrought under the sun is grievous unto me: for all is vanity and vexation of spirit.

uersa sub sole et cuncta uanitatem atque adflictionem spiritus 2:18 rursum detestatus sum omnem industriam meam 2:18 Yea, I hated all my labour which I had taken under

the sun: because I should leave it unto the man that shall be after me.

quae sub sole studiosissime laboraui habiturus heredem post me 2:19 quem ignoro utrum sapiens an stultus futurus sit et 2:19 And who knoweth whether he shall be a wise man or

a fool? yet shall he have rule over all my labour wherein I have laboured, and wherein I have shewed myself wise

dominabitur in laboribus meis quibus desudaui et sollicitus under the sun. This is also vanity.

fui et est quicquam tam uanum 2:20 unde cessaui renun- 2:20 Therefore I went about to cause my heart to despair

of all the labour which I took under the sun.

tiauitque cor meum ultra laborare sub sole 2:21 nam cum 2:21 For there is a man whose labour is in wisdom, and

in knowledge, and in equity; yet to a man that hath not laboured therein shall he leave it for his portion. This also

alius laboret in sapientia et doctrina et sollicitudine homini is vanity and a great evil. otioso quaesita dimittit et hoc ergo uanitas et magnum malum 2:22 quid enim prodeeit homini de uniuerso labore suo 2:22 For what hath man of all his labour, and of the vexa-

tion of his heart, wherein he hath laboured under the sun?

et adflictione spiritus qua sub sole cruciatus est 2:23 cuncti 2:23 For all his days are sorrows, and his travail grief; yea,

his heart taketh not rest in the night. This is also vanity.

dies eius doloribus et aerumnis pleni sunt nec per noctem mente requiescit et haec non uanitas est 2:24 nonne melius 2:24 There is nothing better for a man, than that he should

eat and drink, and that he should make his soul enjoy good in his labour. This also I saw, that it was from the hand of

est comedere et bibere et ostendere animae suae bona de la- God. boribus suis et hoc de manu Dei est 2:25 quis ita uorabit et 2:25 For who can eat, or who else can hasten hereunto,

more than I?

deliciis affluet ut ego 2:26 homini bono in conspectu suo 2:26 For God giveth to a man that is good in his sight wis-

dom, and knowledge, and joy: but to the sinner he giveth travail, to gather and to heap up, that he may give to him that is good before God. This also is vanity and vexation of spirit.

dedit Deus sapientiam et scientiam et laetitiam peccatori autem dedit adflictionem et curam superfluam ut addat et congreget et tradat ei qui placuit Deo sed et hoc uanitas et cassa sollicitudo mentis

3:1 omnia tempus habent et suis spatiis transeunt uniuersa 3:1 To every thing there is a season, and a time to every

purpose under the heaven:

sub caelo 3:2 tempus nascendi et tempus moriendi tempus 3:2 A time to be born, and a time to die; a time to plant,

and a time to pluck up that which is planted;

plantandi et tempus euellendi quod plantatum est 3:3 tem- 3:3 A time to kill, and a time to heal; a time to break down,

and a time to build up;

pus occidendi et tempus sanandi tempus destruendi et tempus aedificandi 3:4 tempus flendi et tempus ridendi tempus 3:4 A time to weep, and a time to laugh; a time to mourn,

and a time to dance;

plangendi et tempus saltandi 3:5 tempus spargendi lapides 3:5 A time to cast away stones, and a time to gather stones

together; a time to embrace, and a time to refrain from embracing;

et tempus colligendi tempus amplexandi et tempus longe fieri a conplexibus 3:6 tempus adquirendi et tempus per- 3:6 A time to get, and a time to lose; a time to keep, and a

time to cast away;

dendi tempus custodiendi et tempus abiciendi 3:7 tempus 3:7 A time to rend, and a time to sew; a time to keep si-

lence, and a time to speak;

scindendi et tempus consuendi tempus tacendi et tempus lo-

Biblia Sacra Vulgata

3:8 A time to love, and a time to hate; a time of war, and a quendi 3:8 tempus dilectionis et tempus odii tempus belli

time of peace.

3:9 What profit hath he that worketh in that wherein he et tempus pacis 3:9 quid habet amplius homo de labore suo

laboureth?

3:10 I have seen the travail, which God hath given to the 3:10 uidi adflictionem quam dedit Deus filiis hominum ut di-

sons of men to be exercised in it. 3:11 He hath made every thing beautiful in his time: also stendantur in ea 3:11 cuncta fecit bona in tempore suo et he hath set the world in their heart, so that no man can find out the work that God maketh from the beginning to the

end. mundum tradidit disputationi eorum ut non inueniat homo 3:12 I know that there is no good in them, but for a man to opus quod operatus est Deus ab initio usque ad finem 3:12

rejoice, and to do good in his life.

et cognoui quod non esset melius nisi laetari et facere bene in

3:13 And also that every man should eat and drink, and uita sua 3:13 omnis enim homo qui comedit et bibit et uidet

enjoy the good of all his labour, it is the gift of God. 3:14 I know that, whatsoever God doeth, it shall be for bonum de labore suo hoc donum Dei est 3:14 didici quod ever: nothing can be put to it, nor any thing taken from it: and God doeth it, that men should fear before him.

omnia opera quae fecit Deus perseuerent in perpetuum non possumus eis quicquam addere nec auferre quae fecit De-

3:15 That which hath been is now; and that which is to be us ut timeatur 3:15 quod factum est ipsum permanet quae

hath already been; and God requireth that which is past.

3:16 And moreover I saw under the sun the place of judg- futura sunt iam fuerunt et Deus instaurat quod abiit 3:16

ment, that wickedness was there; and the place of righteousness, that iniquity was there.

uidi sub sole in loco iudicii impietatem et in loco iustitiae

3:17 I said in mine heart, God shall judge the righteous iniquitatem 3:17 et dixi in corde meo iustum et impium iu-

and the wicked: for there is a time there for every purpose and for every work.

dicabit Deus et tempus omni rei tunc erit 3:18 dixi in corde

3:18 I said in mine heart concerning the estate of the sons of men, that God might manifest them, and that they might see that they themselves are beasts.

meo de filiis hominum ut probaret eos Deus et ostenderet

3:19 For that which befalleth the sons of men befalleth similes esse bestiis 3:19 idcirco unus interitus est homi-

beasts; even one thing befalleth them: as the one dieth, so dieth the other; yea, they have all one breath; so that a man

hath no preeminence above a beast: for all is vanity. nis et iumentorum et aequa utriusque condicio sicut moritur

homo sic et illa moriuntur similiter spirant omnia et nihil ha-

3:20 All go unto one place; all are of the dust, and all turn bet homo iumento amplius cuncta subiacent uanitati 3:20 et

to dust again.

omnia pergunt ad unum locum de terra facta sunt et in ter-

3:21 Who knoweth the spirit of man that goeth upward, ram pariter reuertentur 3:21 quis nouit si spiritus filiorum

and the spirit of the beast that goeth downward to the earth?

Adam ascendat sursum et si spiritus iumentorum descendat

3:22 Wherefore I perceive that there is nothing better, than deorsum 3:22 et deprehendi nihil esse melius quam laetari that a man should rejoice in his own works; for that is his portion: for who shall bring him to see what shall be after

him? hominem in opere suo et hanc esse partem illius quis enim eum adducet ut post se futura cognoscat

4:1 So I returned, and considered all the oppressions that 4:1 uerti me ad alia et uidi calumnias quae sub sole gerunare done under the sun: and behold the tears of such as were oppressed, and they had no comforter; and on the side of their oppressors there was power; but they had no comforter.

tur et lacrimas innocentum et consolatorem neminem nec posse resistere eorum uiolentiae cunctorum auxilio desti-

4:2 Wherefore I praised the dead which are already dead tutos 4:2 et laudaui magis mortuos quam uiuentes 4:3 et more than the living which are yet alive. 4:3 Yea, better is he than both they, which hath not yet been, who hath not seen the evil work that is done under the sun.

feliciorem utroque iudicaui qui necdum natus est nec uidit

4:4 Again, I considered all travail, and every right work, mala quae sub sole fiunt 4:4 rursum contemplatus omnes that for this a man is envied of his neighbour. This is also vanity and vexation of spirit.

Liber Ecclesiastes

labores hominum et industrias animaduerti patere inuidiae

proximi et in hoc ergo uanitas et cura superflua est 4:5 4:5 The fool foldeth his hands together, and eateth his own

flesh.

stultus conplicat manus suas et comedit carnes suas dicens

4:6 melior est pugillus cum requie quam plena utraque ma- 4:6 Better is an handful with quietness, than both the hands

full with travail and vexation of spirit.

nus cum labore et adflictione animi 4:7 considerans rep- 4:7 Then I returned, and I saw vanity under the sun. peri et aliam uanitatem sub sole 4:8 unus est et secundum 4:8 There is one alone, and there is not a second; yea, he

hath neither child nor brother: yet is there no end of all his labour; neither is his eye satisfied with riches; neither saith he, For whom do I labour, and bereave my soul of good? This is also vanity, yea, it is a sore travail.

non habet non filium non fratrem et tamen laborare non cessat nec satiantur oculi eius diuitiis nec recogitat dicens cui laboro et fraudo animam meam bonis in hoc quoque uanitas est et adflictio pessima 4:9 melius ergo est duos simul 4:9 Two are better than one; because they have a good

reward for their labour.

esse quam unum habent enim emolumentum societatis suae 4:10 si unus ceciderit ab altero fulcietur uae soli quia cum 4:10 For if they fall, the one will lift up his fellow: but

woe to him that is alone when he falleth; for he hath not another to help him up.

ruerit non habet subleuantem 4:11 et si dormierint duo fo-

4:11 Again, if two lie together, then they have heat: but how can one be warm alone?

uebuntur mutuo unus quomodo calefiet 4:12 et si quispiam 4:12 And if one prevail against him, two shall withstand

him; and a threefold cord is not quickly broken.

praeualuerit contra unum duo resistent ei funiculus triplex difficile rumpitur 4:13 melior est puer pauper et sapiens re- 4:13 Better is a poor and a wise child than an old and fool-

ish king, who will no more be admonished.

ge sene et stulto qui nescit prouidere in posterum 4:14 quod 4:14 For out of prison he cometh to reign; whereas also he

that is born in his kingdom becometh poor.

et de carcere catenisque interdum quis egrediatur ad regnum et alius natus in regno inopia consumatur 4:15 uidi cunctos 4:15 I considered all the living which walk under the sun,

with the second child that shall stand up in his stead.

uiuentes qui ambulant sub sole cum adulescente secundo qui consurgit pro eo 4:16 infinitus numerus est populi omnium 4:16 There is no end of all the people, even of all that

have been before them: they also that come after shall not rejoice in him. Surely this also is vanity and vexation of

qui fuerunt ante eum et qui postea futuri sunt non laetabun- spirit. tur in eo sed et hoc uanitas et adflictio spiritus 4:17 custodi 4:17 Keep thy foot when thou goest to the house of God,

and be more ready to hear, than to give the sacrifice of fools: for they consider not that they do evil.

pedem tuum ingrediens domum Dei multo enim melior est oboedientia quam stultorum uictimae qui nesciunt quid faciant mali

5:1 ne temere quid loquaris neque cor tuum sit uelox ad 5:1 Be not rash with thy mouth, and let not thine heart be hasty to utter any thing before God: for God is in heaven, and thou upon earth: therefore let thy words be few.

proferendum sermonem coram Deo Deus enim in caelo et tu super terram idcirco sint pauci sermones tui 5:2 multas cu- 5:2 For a dream cometh through the multitude of business;

and a fool’s voice is known by multitude of words.

ras sequuntur somnia et in multis sermonibus inuenitur stultitia 5:3 si quid uouisti Deo ne moreris reddere displicet 5:3 When thou vowest a vow unto God, defer not to pay it;

for he hath no pleasure in fools: pay that which thou hast vowed.

enim ei infidelis et stulta promissio sed quodcumque uoueris

redde 5:4 multoque melius est non uouere quam post uo- 5:4 Better is it that thou shouldest not vow, than that thou shouldest vow and not pay.

Biblia Sacra Vulgata

5:5 Suffer not thy mouth to cause thy flesh to sin; neither tum promissa non conplere 5:5 ne dederis os tuum ut pecsay thou before the angel, that it was an error: wherefore should God be angry at thy voice, and destroy the work of

thine hands? care faciat carnem tuam neque dicas coram angelo non est

prouidentia ne forte iratus Deus super sermone tuo dissipet

5:6 For in the multitude of dreams and many words there cuncta opera manuum tuarum 5:6 ubi multa sunt somnia are also divers vanities: but fear thou God.

plurimae uanitates et sermones innumeri tu uero Deum time

5:7 If thou seest the oppression of the poor, and violent 5:7 si uideris calumnias egenorum et uiolenta iudicia et sub-

perverting of judgment and justice in a province, marvel not at the matter: for he that is higher than the highest

regardeth; and there be higher than they. uerti iustitiam in prouincia non mireris super hoc negotio

quia excelso alius excelsior est et super hos quoque eminen-

5:8 Moreover the profit of the earth is for all: the king tiores sunt alii 5:8 et insuper uniuersae terrae rex imperat

himself is served by the field. 5:9 He that loveth silver shall not be satisfied with silver; seruienti 5:9 auarus non implebitur pecunia et qui amat dinor he that loveth abundance with increase: this is also vanity.

uitias fructus non capiet ex eis et hoc ergo uanitas 5:10 ubi

5:10 When goods increase, they are increased that eat them: and what good is there to the owners thereof, saving the beholding of them with their eyes?

multae sunt opes multi et qui comedant eas et quid prodeet

5:11 The sleep of a labouring man is sweet, whether he eat possessori nisi quod cernit diuitias oculis suis 5:11 dulcis

little or much: but the abundance of the rich will not suffer him to sleep.

est somnus operanti siue parum siue multum comedat satu-

5:12 There is a sore evil which I have seen under the sun, ritas autem diuitis non sinit dormire eum 5:12 est et alia

namely, riches kept for the owners thereof to their hurt.

infirmitas pessima quam uidi sub sole diuitiae conseruatae

5:13 But those riches perish by evil travail: and he beget- in malum domini sui 5:13 pereunt enim in adflictione pesteth a son, and there is nothing in his hand.

5:14 As he came forth of his mother’s womb, naked shall sima generauit filium qui in summa egestate erit 5:14 sicut

he return to go as he came, and shall take nothing of his labour, which he may carry away in his hand.

egressus est nudus de utero matris suae sic reuertetur et nihil

5:15 And this also is a sore evil, that in all points as he auferet secum de labore suo 5:15 miserabilis prorsus infir-

came, so shall he go: and what profit hath he that hath laboured for the wind?

mitas quomodo uenit sic reuertetur quid ergo prodeet ei quod

5:16 All his days also he eateth in darkness, and he hath laborauit in uentum 5:16 cunctis diebus uitae suae come-

much sorrow and wrath with his sickness.

dit in tenebris et in curis multis et in aerumna atque tristitia

5:17 Behold that which I have seen: it is good and comely 5:17 hoc itaque mihi uisum est bonum ut comedat quis et bi-

for one to eat and to drink, and to enjoy the good of all his labour that he taketh under the sun all the days of his life,

which God giveth him: for it is his portion. bat et fruatur laetitia ex labore suo quod laborauit ipse sub

sole numerum dierum uitae suae quos dedit ei Deus et haec

5:18 Every man also to whom God hath given riches and est pars illius 5:18 et omni homini cui dedit Deus diuitias

wealth, and hath given him power to eat thereof, and to take his portion, and to rejoice in his labour; this is the gift

of God. atque substantiam potestatemque ei tribuit ut comedat ex eis

et fruatur parte sua et laetetur de labore suo hoc est donum

5:19 For he shall not much remember the days of his life; Dei 5:19 non enim satis recordabitur dierum uitae suae eo

because God answereth him in the joy of his heart.

quod Deus occupet deliciis cor eius

6:1 There is an evil which I have seen under the sun, and 6:1 est et aliud malum quod uidi sub sole et quidem freit is common among men:

6:2 A man to whom God hath given riches, wealth, and quens apud homines 6:2 uir cui dedit Deus diuitias et sub-

honour, so that he wanteth nothing for his soul of all that he desireth, yet God giveth him not power to eat thereof, but a stranger eateth it: this is vanity, and it is an evil disease.

Liber Ecclesiastes

stantiam et honorem et nihil deest animae eius ex omnibus quae desiderat nec tribuit ei potestatem Deus ut comedat ex eo sed homo extraneus uorabit illud hoc uanitas et magna miseria est 6:3 si genuerit quispiam centum et uixerit mul- 6:3 If a man beget an hundred children, and live many

years, so that the days of his years be many, and his soul be not filled with good, and also that he have no burial; I

tos annos et plures dies aetatis habuerit et anima illius non say, that an untimely birth is better than he. utatur bonis substantiae suae sepulturaque careat de hoc ego pronuntio quod melior illo sit abortiuus 6:4 frustra enim 6:4 For he cometh in with vanity, and departeth in dark-

ness, and his name shall be covered with darkness.

uenit et pergit ad tenebras et obliuione delebitur nomen eius 6:5 non uidit solem neque cognouit distantiam boni et mali 6:5 Moreover he hath not seen the sun, nor known any

thing: this hath more rest than the other.

6:6 etiam si duobus milibus annis uixerit et non fuerit per- 6:6 Yea, though he live a thousand years twice told, yet

hath he seen no good: do not all go to one place?

fruitus bonis nonne ad unum locum properant omnia 6:7 6:7 All the labour of man is for his mouth, and yet the

appetite is not filled.

omnis labor hominis in ore eius sed anima illius non impletur 6:8 quid habet amplius sapiens ab stulto et quid pauper 6:8 For what hath the wise more than the fool? what hath

the poor, that knoweth to walk before the living?

nisi ut pergat illuc ubi est uita 6:9 melius est uidere quod 6:9 Better is the sight of the eyes than the wandering of

the desire: this is also vanity and vexation of spirit.

cupias quam desiderare quod nescias sed et hoc uanitas est et praesumptio spiritus 6:10 qui futurus est iam uocatum 6:10 That which hath been is named already, and it is

known that it is man: neither may he contend with him that is mightier than he.

est nomen eius et scitur quod homo sit et non possit contra fortiorem se in iudicio contendere 6:11 uerba sunt plurima 6:11 Seeing there be many things that increase vanity,

what is man the better?

multa in disputando habentia uanitatem

7:1 quid necesse est homini maiora se quaerere cum igno- 7:1 For who knoweth what is good for man in this life, all

the days of his vain life which he spendeth as a shadow? for who can tell a man what shall be after him under the

ret quid conducat sibi in uita sua numero dierum peregrina- sun? tionis suae et tempore quo uelut umbra praeterit aut quis ei poterit indicare quid post eum futurum sub sole sit 7:2 me- 7:2 A good name is better than precious ointment; and the

day of death than the day of one’s birth.

lius est nomen bonum quam unguenta pretiosa et dies mortis die natiuitatis 7:3 melius est ire ad domum luctus quam ad 7:3 It is better to go to the house of mourning, than to go

to the house of feasting: for that is the end of all men; and the living will lay it to his heart.

domum conuiuii in illa enim finis cunctorum admonetur hominum et uiuens cogitat quid futurum sit 7:4 melior est ira 7:4 Sorrow is better than laughter: for by the sadness of

the countenance the heart is made better.

risu quia per tristitiam uultus corrigitur animus delinquentis 7:5 cor sapientium ubi tristitia est et cor stultorum ubi lae- 7:5 The heart of the wise is in the house of mourning; but

the heart of fools is in the house of mirth.

titia 7:6 melius est a sapiente corripi quam stultorum ad- 7:6 It is better to hear the rebuke of the wise, than for a

man to hear the song of fools.

ulatione decipi 7:7 quia sicut sonitus spinarum ardentium 7:7 For as the crackling of thorns under a pot, so is the

laughter of the fool: this also is vanity.

sub olla sic risus stulti sed et hoc uanitas 7:8 calumnia con- 7:8 Surely oppression maketh a wise man mad; and a gift

destroyeth the heart.

turbat sapientem et perdet robur cordis illius 7:9 melior est 7:9 Better is the end of a thing than the beginning thereof:

and the patient in spirit is better than the proud in spirit.

Biblia Sacra Vulgata

finis orationis quam principium melior est patiens arrogante

7:10 Be not hasty in thy spirit to be angry: for anger resteth 7:10 ne uelox sis ad irascendum quia ira in sinu stulti requies-

in the bosom of fools.

7:11 Say not thou, What is the cause that the former days cit 7:11 ne dicas quid putas causae est quod priora tempora

were better than these? for thou dost not enquire wisely concerning this.

meliora fuere quam nunc sunt stulta est enim huiuscemodi

7:12 Wisdom is good with an inheritance: and by it there interrogatio 7:12 utilior est sapientia cum diuitiis et magis

is profit to them that see the sun.

7:13 For wisdom is a defence, and money is a defence: but prodeet uidentibus solem 7:13 sicut enim protegit sapientia

the excellency of knowledge is, that wisdom giveth life to them that have it.

sic protegit pecunia hoc autem plus habet eruditio et sapien-

7:14 Consider the work of God: for who can make that tia quod uitam tribuunt possessori suo 7:14 considera opera

straight, which he hath made crooked?

7:15 In the day of prosperity be joyful, but in the day of Dei quod nemo possit corrigere quem ille despexerit 7:15

adversity consider: God also hath set the one over against the other, to the end that man should find nothing after

him. in die bona fruere bonis et malam diem praecaue sicut enim

hanc sic et illam fecit Deus ut non inueniat homo contra eum

7:16 All things have I seen in the days of my vanity: there iustas querimonias 7:16 haec quoque uidi in diebus uani-

is a just man that perisheth in his righteousness, and there is a wicked man that prolongeth his life in his wickedness.

tatis meae iustus perit in iustitia sua et impius multo uiuit

7:17 Be not righteous over much; neither make thyself tempore in malitia sua 7:17 noli esse iustus multum neque

over wise: why shouldest thou destroy thyself ?

7:18 Be not over much wicked, neither be thou foolish: plus sapias quam necesse est ne obstupescas 7:18 ne im-

why shouldest thou die before thy time?

pie agas multum et noli esse stultus ne moriaris in tempore

7:19 It is good that thou shouldest take hold of this; yea, non tuo 7:19 bonum est te sustentare iustum sed et ab illo also from this withdraw not thine hand: for he that feareth God shall come forth of them all.

ne subtrahas manum tuam quia qui Deum timet nihil negle-

7:20 Wisdom strengtheneth the wise more than ten mighty git 7:20 sapientia confortabit sapientem super decem prin-

men which are in the city.

7:21 For there is not a just man upon earth, that doeth cipes ciuitatis 7:21 non est enim homo iustus in terra qui

good, and sinneth not. 7:22 Also take no heed unto all words that are spoken; lest faciat bonum et non peccet 7:22 sed et cunctis sermonibus thou hear thy servant curse thee:

qui dicuntur ne accommodes cor tuum ne forte audias ser-

7:23 For oftentimes also thine own heart knoweth that thou uum tuum maledicentem tibi 7:23 scit enim tua conscientia thyself likewise hast cursed others.

7:24 All this have I proved by wisdom: I said, I will be quia et tu crebro maledixisti aliis 7:24 cuncta temptaui in

wise; but it was far from me.

sapientia dixi sapiens efficiar et ipsa longius recessit a me

7:25 That which is far off, and exceeding deep, who can 7:25 multo magis quam erat et alta profunditas quis inueniet

find it out?

7:26 I applied mine heart to know, and to search, and to eam 7:26 lustraui uniuersa animo meo ut scirem et consi-

seek out wisdom, and the reason of things, and to know the wickedness of folly, even of foolishness and madness:

derarem et quaererem sapientiam et rationem et ut cognos-

7:27 And I find more bitter than death the woman, whose cerem impietatem stulti et errorem inprudentium 7:27 et in-

heart is snares and nets, and her hands as bands: whoso pleaseth God shall escape from her; but the sinner shall be

taken by her. ueni amariorem morte mulierem quae laqueus uenatorum est

et sagena cor eius uincula sunt manus illius qui placet Deo

7:28 Behold, this have I found, saith the preacher, count- effugiet eam qui autem peccator est capietur ab illa 7:28 ing one by one, to find out the account:

ecce hoc inueni dicit Ecclesiastes unum et alterum ut inue-

Liber Ecclesiastes

nirem rationem 7:29 quam adhuc quaerit anima mea et non 7:29 Which yet my soul seeketh, but I find not: one man

among a thousand have I found; but a woman among all those have I not found.

inueni uirum de mille unum repperi mulierem ex omnibus non inueni 7:30 solummodo hoc inueni quod fecerit Deus 7:30 Lo, this only have I found, that God hath made man

upright; but they have sought out many inventions.

hominem rectum et ipse se infinitis miscuerit quaestionibus quis talis ut sapiens est et quis cognouit solutionem uerbi

8:1 sapientia hominis lucet in uultu eius et potentissimus 8:1 Who is as the wise man? and who knoweth the inter-

pretation of a thing? a man’s wisdom maketh his face to shine, and the boldness of his face shall be changed.

faciem illius commutauit 8:2 ego os regis obseruo et prae-

8:2 I counsel thee to keep the king’s commandment, and that in regard of the oath of God.

cepta iuramenti Dei 8:3 ne festines recedere a facie eius _{8:3 Be not hasty to go out of his sight: stand not in an evil}

thing; for he doeth whatsoever pleaseth him.

neque permaneas in opere malo quia omne quod uoluerit faciet 8:4 et sermo illius potestate plenus est nec dicere ei 8:4 Where the word of a king is, there is power: and who

may say unto him, What doest thou?

quisquam potest quare ita facis 8:5 qui custodit praeceptum 8:5 Whoso keepeth the commandment shall feel no evil

thing: and a wise man’s heart discerneth both time and judgment.

non experietur quicquam mali tempus et responsionem cor sapientis intellegit 8:6 omni negotio tempus est et oportuni- 8:6 Because to every purpose there is time and judgment,

therefore the misery of man is great upon him.

tas et multa hominis adflictio 8:7 quia ignorat praeterita et 8:7 For he knoweth not that which shall be: for who can

tell him when it shall be?

uentura nullo scire potest nuntio 8:8 non est in hominis di- 8:8 There is no man that hath power over the spirit to retain

the spirit; neither hath he power in the day of death: and there is no discharge in that war; neither shall wickedness

cione prohibere spiritum nec habet potestatem in die mortis deliver those that are given to it. nec sinitur quiescere ingruente bello neque saluabit impie-

tas impium 8:9 omnia haec consideraui et dedi cor meum 8:9 All this have I seen, and applied my heart unto every work that is done under the sun: there is a time wherein one man ruleth over another to his own hurt.

in cunctis operibus quae fiunt sub sole interdum dominatur homo homini in malum suum 8:10 uidi impios sepultos qui 8:10 And so I saw the wicked buried, who had come and

gone from the place of the holy, and they were forgotten in the city where they had so done: this is also vanity.

etiam cum aduiuerent in loco sancto erant et laudabantur in ciuitate quasi iustorum operum sed et hoc uanitas est 8:11 8:11 Because sentence against an evil work is not executed

speedily, therefore the heart of the sons of men is fully set in them to do evil.

etenim quia non profertur cito contra malos sententia absque ullo timore filii hominum perpetrant mala 8:12 attamen ex 8:12 Though a sinner do evil an hundred times, and his

days be prolonged, yet surely I know that it shall be well with them that fear God, which fear before him:

eo quod peccator centies facit malum et per patientiam sustentatur ego cognoui quod erit bonum timentibus Deum qui

uerentur faciem eius 8:13 non sit bonum impio nec prolon- 8:13 But it shall not be well with the wicked, neither shall he prolong his days, which are as a shadow; because he feareth not before God.

gentur dies eius sed quasi umbra transeant qui non timent faciem Dei 8:14 est et alia uanitas quae fit super terram sunt 8:14 There is a vanity which is done upon the earth; that

there be just men, unto whom it happeneth according to the work of the wicked; again, there be wicked men, to whom it happeneth according to the work of the righteous: I said that this also is vanity.

iusti quibus multa proueniunt quasi opera egerint impiorum et sunt impii qui ita securi sunt quasi iustorum facta habeant sed et hoc uanissimum iudico 8:15 laudaui igitur laetitiam 8:15 Then I commended mirth, because a man hath no

better thing under the sun, than to eat, and to drink, and to be merry: for that shall abide with him of his labour the

quod non esset homini bonum sub sole nisi quod comede- days of his life, which God giveth him under the sun.

Biblia Sacra Vulgata

ret et biberet atque gauderet et hoc solum secum auferret de labore suo in diebus uitae quos dedit ei Deus sub sole

8:16 When I applied mine heart to know wisdom, and to 8:16 et adposui cor meum ut scirem sapientiam et intellege-

see the business that is done upon the earth: (for also there is that neither day nor night seeth sleep with his eyes:)

rem distentionem quae uersatur in terra est homo qui diebus

8:17 Then I beheld all the work of God, that a man can- ac noctibus somnum oculis non capit 8:17 et intellexi quod

not find out the work that is done under the sun: because though a man labour to seek it out, yet he shall not find it; yea farther; though a wise man think to know it, yet shall he not be able to find it.

omnium operum Dei nullam possit homo inuenire rationem eorum quae fiunt sub sole et quanto plus laborauerit ad quaerendum tanto minus inueniat etiam si dixerit sapiens se nosse non poterit repperire

9:1 For all this I considered in my heart even to declare all 9:1 omnia haec tractaui in corde meo ut curiose intellegethis, that the righteous, and the wise, and their works, are in the hand of God: no man knoweth either love or hatred

by all that is before them. rem sunt iusti atque sapientes et opera eorum in manu Dei

9:2 All things come alike to all: there is one event to the et tamen nescit homo utrum amore an odio dignus sit 9:2 righteous, and to the wicked; to the good and to the clean, and to the unclean; to him that sacrificeth, and to him that sacrificeth not: as is the good, so is the sinner; and he that sweareth, as he that feareth an oath.

sed omnia in futuro seruantur incerta eo quod uniuersa aeque eueniant iusto et impio bono et malo mundo et inmundo immolanti uictimas et sacrificia contemnenti sicut bonus sic

9:3 This is an evil among all things that are done under the et peccator ut periurus ita et ille qui uerum deierat 9:3 hoc

sun, that there is one event unto all: yea, also the heart of the sons of men is full of evil, and madness is in their heart

while they live, and after that they go to the dead. est pessimum inter omnia quae sub sole fiunt quia eadem cunctis eueniunt unde et corda filiorum hominum implentur malitia et contemptu in uita sua et post haec ad inferos dedu-

9:4 For to him that is joined to all the living there is hope: centur 9:4 nemo est qui semper uiuat et qui huius rei habeat for a living dog is better than a dead lion.

9:5 For the living know that they shall die: but the dead fiduciam melior est canis uiuens leone mortuo 9:5 uiuen-

know not any thing, neither have they any more a reward; for the memory of them is forgotten.

tes enim sciunt se esse morituros mortui uero nihil nouerunt amplius nec habent ultra mercedem quia obliuioni tra-

9:6 Also their love, and their hatred, and their envy, is now dita est memoria eorum 9:6 amor quoque et odium et inperished; neither have they any more a portion for ever in any thing that is done under the sun.

uidia simul perierunt nec habent partem in hoc saeculo et

9:7 Go thy way, eat thy bread with joy, and drink thy wine in opere quod sub sole geritur 9:7 uade ergo et comede in with a merry heart; for God now accepteth thy works.

laetitia panem tuum et bibe cum gaudio uinum tuum quia

9:8 Let thy garments be always white; and let thy head Deo placent opera tua 9:8 omni tempore sint uestimenta tua

lack no ointment.

9:9 Live joyfully with the wife whom thou lovest all the candida et oleum de capite tuo non deficiat 9:9 perfruere

days of the life of thy vanity, which he hath given thee under the sun, all the days of thy vanity: for that is thy portion in this life, and in thy labour which thou takest under the sun.

uita cum uxore quam diligis cunctis diebus uitae instabilitatis tuae qui dati sunt tibi sub sole omni tempore uanitatis tuae haec est enim pars in uita et in labore tuo quod laboras

9:10 Whatsoever thy hand findeth to do, do it with thy sub sole 9:10 quodcumque potest manus tua facere instan-

might; for there is no work, nor device, nor knowledge, nor wisdom, in the grave, whither thou goest.

Liber Ecclesiastes

ter operare quia nec opus nec ratio nec scientia nec sapientia

erunt apud inferos quo tu properas 9:11 uerti me alio ui- 9:11 I returned, and saw under the sun, that the race is not to the swift, nor the battle to the strong, neither yet bread to the wise, nor yet riches to men of understanding, nor yet favour to men of skill; but time and chance happeneth to them all.

dique sub sole nec uelocium esse cursum nec fortium bellum nec sapientium panem nec doctorum diuitias nec artificum gratiam sed tempus casumque in omnibus 9:12 nescit 9:12 For man also knoweth not his time: as the fishes that

are taken in an evil net, and as the birds that are caught in the snare; so are the sons of men snared in an evil time,

homo finem suum sed sicut pisces capiuntur hamo et sic- when it falleth suddenly upon them. ut aues conprehenduntur laqueo sic capiuntur homines tempore malo cum eis extemplo superuenerit 9:13 hanc quoque 9:13 This wisdom have I seen also under the sun, and it

seemed great unto me:

uidi sub sole sapientiam et probaui maximam 9:14 ciuitas 9:14 There was a little city, and few men within it; and

there came a great king against it, and besieged it, and built great bulwarks against it:

parua et pauci in ea uiri uenit contra eam rex magnus et uallauit eam extruxitque munitiones per gyrum et perfecta est obsidio 9:15 inuentusque in ea uir pauper et sapiens libera- 9:15 Now there was found in it a poor wise man, and he

by his wisdom delivered the city; yet no man remembered that same poor man.

uit urbem per sapientiam suam et nullus deinceps recordatus est hominis illius pauperis 9:16 et dicebam ego meliorem 9:16 Then said I, Wisdom is better than strength: never-

theless the poor man’s wisdom is despised, and his words are not heard.

esse sapientiam fortitudine quomodo ergo sapientia pauperis contempta est et uerba eius non sunt audita 9:17 uerba 9:17 The words of wise men are heard in quiet more than

the cry of him that ruleth among fools.

sapientium audiuntur in silentio plus quam clamor principis inter stultos 9:18 melior est sapientia quam arma bellica et 9:18 Wisdom is better than weapons of war: but one sinner

destroyeth much good.

qui in uno peccauerit multa bona perdet

10:1 muscae morientes perdunt suauitatem unguenti pre- 10:1 Dead flies cause the ointment of the apothecary to

send forth a stinking savour: so doth a little folly him that is in reputation for wisdom and honour.

tiosior est sapientia et gloria parua ad tempus stultitia 10:2

10:2 A wise man’s heart is at his right hand; but a fool’s heart at his left.

cor sapientis in dextera eius et cor stulti in sinistra illius 10:3 sed et in uia stultus ambulans cum ipse insipiens sit 10:3 Yea also, when he that is a fool walketh by the way,

his wisdom faileth him, and he saith to every one that he is a fool.

omnes stultos aestimat 10:4 si spiritus potestatem haben-

10:4 If the spirit of the ruler rise up against thee, leave not thy place; for yielding pacifieth great offences.

tis ascenderit super te locum tuum ne dimiseris quia curatio cessare faciet peccata maxima 10:5 est malum quod uidi 10:5 There is an evil which I have seen under the sun, as

an error which proceedeth from the ruler:

sub sole quasi per errorem egrediens a facie principis 10:6 10:6 Folly is set in great dignity, and the rich sit in low

place.

positum stultum in dignitate sublimi et diuites sedere deorsum 10:7 uidi seruos in equis et principes ambulantes quasi 10:7 I have seen servants upon horses, and princes walking

as servants upon the earth.

seruos super terram 10:8 qui fodit foueam incidet in eam et 10:8 He that diggeth a pit shall fall into it; and whoso

breaketh an hedge, a serpent shall bite him.

qui dissipat sepem mordebit eum coluber 10:9 qui transfert 10:9 Whoso removeth stones shall be hurt therewith; and

he that cleaveth wood shall be endangered thereby.

lapides adfligetur in eis et qui scindit ligna uulnerabitur ab eis 10:10 si retunsum fuerit ferrum et hoc non ut prius sed 10:10 If the iron be blunt, and he do not whet the edge,

then must he put to more strength: but wisdom is profitable to direct.

Biblia Sacra Vulgata

hebetatum erit multo labore exacuatur et post industriam se-

10:11 Surely the serpent will bite without enchantment; quitur sapientia 10:11 si mordeat serpens in silentio nihil eo

and a babbler is no better.

10:12 The words of a wise man’s mouth are gracious; but minus habet qui occulte detrahit 10:12 uerba oris sapientis

the lips of a fool will swallow up himself.

10:13 The beginning of the words of his mouth is foolish- gratia et labia insipientis praecipitabunt eum 10:13 initium

ness: and the end of his talk is mischievous madness.

uerborum eius stultitia et nouissimum oris illius error pessi-

10:14 A fool also is full of words: a man cannot tell what mus 10:14 stultus uerba multiplicat ignorat homo quid ante

shall be; and what shall be after him, who can tell him?

se fuerit et quod post futurum est quis illi poterit indicare

10:15 The labour of the foolish wearieth every one of 10:15 labor stultorum adfliget eos qui nesciunt in urbem per-

them, because he knoweth not how to go to the city.

10:16 Woe to thee, O land, when thy king is a child, and gere 10:16 uae tibi terra cuius rex est puer et cuius princi-

thy princes eat in the morning!

10:17 Blessed art thou, O land, when thy king is the son pes mane comedunt 10:17 beata terra cuius rex nobilis est et

of nobles, and thy princes eat in due season, for strength, and not for drunkenness!

cuius principes uescuntur in tempore suo ad reficiendum et

10:18 By much slothfulness the building decayeth; and non ad luxuriam 10:18 in pigritiis humiliabitur contignatio

through idleness of the hands the house droppeth through.

10:19 A feast is made for laughter, and wine maketh et in infirmitate manuum perstillabit domus 10:19 in risu

merry: but money answereth all things.

faciunt panem ac uinum ut epulentur uiuentes et pecuniae

10:20 Curse not the king, no not in thy thought; and curse oboedient omnia 10:20 in cogitatione tua regi ne detrahas

not the rich in thy bedchamber: for a bird of the air shall carry the voice, and that which hath wings shall tell the

matter. et in secreto cubiculi tui ne maledixeris diuiti quia auis caeli portabit uocem tuam et qui habet pinnas adnuntiabit sententiam

11:1 Cast thy bread upon the waters: for thou shalt find it 11:1 mitte panem tuum super transeuntes aquas quia post after many days.

11:2 Give a portion to seven, and also to eight; for thou multa tempora inuenies illum 11:2 da partem septem nec-

knowest not what evil shall be upon the earth.

non et octo quia ignoras quid futurum sit mali super terram

11:3 If the clouds be full of rain, they empty themselves 11:3 si repletae fuerint nubes imbrem super terram effundent

upon the earth: and if the tree fall toward the south, or toward the north, in the place where the tree falleth, there

it shall be. si ceciderit lignum ad austrum aut ad aquilonem in quocum-11:4 He that observeth the wind shall not sow; and he that que loco ceciderit ibi erit 11:4 qui obseruat uentum non se-

regardeth the clouds shall not reap. 11:5 As thou knowest not what is the way of the spirit, nor minat et qui considerat nubes numquam metet 11:5 quohow the bones do grow in the womb of her that is with child: even so thou knowest not the works of God who

maketh all. modo ignoras quae sit uia spiritus et qua ratione conpingan-

tur ossa in uentre praegnatis sic nescis opera Dei qui fab-

11:6 In the morning sow thy seed, and in the evening with- ricator est omnium 11:6 mane semina sementem tuam et hold not thine hand: for thou knowest not whether shall prosper, either this or that, or whether they both shall be

alike good. uespere ne cesset manus tua quia nescis quid magis oria-

11:7 Truly the light is sweet, and a pleasant thing it is for tur hoc an illud et si utrumque simul melius erit 11:7 dulce the eyes to behold the sun:

11:8 But if a man live many years, and rejoice in them all; lumen et delectabile est oculis uidere solem 11:8 si annis

yet let him remember the days of darkness; for they shall be many. All that cometh is vanity.

multis uixerit homo et in omnibus his laetatus fuerit meminisse debet tenebrosi temporis et dierum multorum qui cum

Liber Ecclesiastes

uenerint uanitatis arguentur praeterita 11:9 laetare ergo iu- 11:9 Rejoice, O young man, in thy youth; and let thy heart

cheer thee in the days of thy youth, and walk in the ways of thine heart, and in the sight of thine eyes: but know thou,

uenis in adulescentia tua et in bono sit cor tuum in diebus that for all these things God will bring thee into judgment.

iuuentutis tuae et ambula in uiis cordis tui et in intuitu oculorum tuorum et scito quod pro omnibus his adducet te Deus in iudicium 11:10 aufer iram a corde tuo et amoue malitiam 11:10 Therefore remove sorrow from thy heart, and put

away evil from thy flesh: for childhood and youth are vanity.

a carne tua adulescentia enim et uoluptas uana sunt

12:1 memento creatoris tui in diebus iuuentutis tuae an- 12:1 Remember now thy Creator in the days of thy youth,

while the evil days come not, nor the years draw nigh, when thou shalt say, I have no pleasure in them;

tequam ueniat tempus adflictionis et adpropinquent anni de quibus dicas non mihi placent 12:2 antequam tenebrescat 12:2 While the sun, or the light, or the moon, or the stars,

be not darkened, nor the clouds return after the rain:

sol et lumen et luna et stellae et reuertantur nubes post plu-

uiam 12:3 quando commouebuntur custodes domus et nu- 12:3 In the day when the keepers of the house shall tremble, and the strong men shall bow themselves, and the grinders cease because they are few, and those that look

tabuntur uiri fortissimi et otiosae erunt molentes inminuto out of the windows be darkened, numero et tenebrescent uidentes per foramina 12:4 et clau- 12:4 And the doors shall be shut in the streets, when the

sound of the grinding is low, and he shall rise up at the voice of the bird, and all the daughters of musick shall be

dent ostia in platea in humilitate uocis molentis et consur- brought low; gent ad uocem uolucris et obsurdescent omnes filiae carminis 12:5 excelsa quoque timebunt et formidabunt in uia flo- 12:5 Also when they shall be afraid of that which is high,

and fears shall be in the way, and the almond tree shall flourish, and the grasshopper shall be a burden, and desire shall fail: because man goeth to his long home, and the mourners go about the streets:

rebit amigdalum inpinguabitur lucusta et dissipabitur capparis quoniam ibit homo in domum aeternitatis suae et circumibunt in platea plangentes 12:6 antequam rumpatur fu- 12:6 Or ever the silver cord be loosed, or the golden bowl

be broken, or the pitcher be broken at the fountain, or the wheel broken at the cistern.

nis argenteus et recurrat uitta aurea et conteratur hydria super fontem et confringatur rota super cisternam 12:7 et re- 12:7 Then shall the dust return to the earth as it was: and

the spirit shall return unto God who gave it.

uertatur puluis in terram suam unde erat et spiritus redeet

ad Deum qui dedit illum 12:8 uanitas uanitatum dixit Ec- 12:8 Vanity of vanities, saith the preacher; all is vanity.

clesiastes omnia uanitas 12:9 cumque esset sapientissimus 12:9 And moreover, because the preacher was wise, he

still taught the people knowledge; yea, he gave good heed, and sought out, and set in order many proverbs.

Ecclesiastes docuit populum et enarrauit quae fecerit et inuestigans conposuit parabolas multas 12:10 quaesiuit uerba 12:10 The preacher sought to find out acceptable words:

and that which was written was upright, even words of truth.

utilia et conscripsit sermones rectissimos ac ueritate plenos 12:11 uerba sapientium sicut stimuli et quasi claui in altum 12:11 The words of the wise are as goads, and as nails fas-

tened by the masters of assemblies, which are given from one shepherd.

defixi quae per magistrorum concilium data sunt a pastore uno 12:12 his amplius fili mi ne requiras faciendi plures lib- 12:12 And further, by these, my son, be admonished: of

making many books there is no end; and much study is a weariness of the flesh.

ros nullus est finis frequensque meditatio carnis adflictio est

12:13 finem loquendi omnes pariter audiamus Deum time et 12:13 Let us hear the conclusion of the whole matter: Fear God, and keep his commandments: for this is the whole duty of man.

mandata eius obserua hoc est enim omnis homo 12:14 et

12:14 For God shall bring every work into judgment, with every secret thing, whether it be good, or whether it be evil.

Biblia Sacra Vulgata

cuncta quae fiunt adducet Deus in iudicium pro omni errato siue bonum siue malum sit

Canticum Canticorum

1:1 osculetur me osculo oris sui quia meliora sunt ubera 1:1 Let him kiss me with the kisses of his mouth: for thy

love is better than wine.

tua uino 1:2 fraglantia unguentis optimis oleum effusum 1:2 Because of the savour of thy good ointments thy name

is as ointment poured forth, therefore do the virgins love thee.

nomen tuum ideo adulescentulae dilexerunt te 1:3 trahe me

1:3 Draw me, we will run after thee: the king hath brought me into his chambers: we will be glad and rejoice in thee, we will remember thy love more than wine: the upright love thee.

post te curremus introduxit me rex in cellaria sua exultabimus et laetabimur in te memores uberum tuorum super uinum recti diligunt te 1:4 nigra sum sed formonsa filiae Hie- 1:4 I am black, but comely, O ye daughters of Jerusalem,

as the tents of Kedar, as the curtains of Solomon.

rusalem sicut tabernacula Cedar sicut pelles Salomonis 1:5 1:5 Look not upon me, because I am black, because the

sun hath looked upon me: my mother’s children were angry with me; they made me the keeper of the vineyards;

nolite me considerare quod fusca sim quia decolorauit me but mine own vineyard have I not kept. sol filii matris meae pugnauerunt contra me posuerunt me custodem in uineis uineam meam non custodiui 1:6 indica 1:6 Tell me, O thou whom my soul loveth, where thou

feedest, where thou makest thy flock to rest at noon: for why should I be as one that turneth aside by the flocks of

mihi quem diligit anima mea ubi pascas ubi cubes in meridie thy companions? ne uagari incipiam per greges sodalium tuorum 1:7 si igno- 1:7 If thou know not, O thou fairest among women, go thy

way forth by the footsteps of the flock, and feed thy kids beside the shepherds’ tents.

ras te o pulchra inter mulieres egredere et abi post uestigia gregum et pasce hedos tuos iuxta tabernacula pastorum 1:8 1:8 I have compared thee, O my love, to a company of

horses in Pharaoh’s chariots.

equitatui meo in curribus Pharaonis adsimilaui te amica mea 1:9 pulchrae sunt genae tuae sicut turturis collum tuum sic- 1:9 Thy cheeks are comely with rows of jewels, thy neck

with chains of gold.

ut monilia 1:10 murenulas aureas faciemus tibi uermicula- 1:10 We will make thee borders of gold with studs of sil-

ver.

tas argento 1:11 dum esset rex in accubitu suo nardus mea 1:11 While the king sitteth at his table, my spikenard

sendeth forth the smell thereof.

dedit odorem suum 1:12 fasciculus murrae dilectus meus 1:12 A bundle of myrrh is my well-beloved unto me; he

shall lie all night betwixt my breasts.

mihi inter ubera mea commorabitur 1:13 botrus cypri di- 1:13 My beloved is unto me as a cluster of camphire in the

vineyards of Engedi.

lectus meus mihi in uineis Engaddi 1:14 ecce tu pulchra es 1:14 Behold, thou art fair, my love; behold, thou art fair;

thou hast doves’ eyes.

amica mea ecce tu pulchra oculi tui columbarum 1:15 ecce 1:15 Behold, thou art fair, my beloved, yea, pleasant: also

our bed is green.

tu pulcher es dilecte mi et decorus lectulus noster floridus 1:16 tigna domorum nostrarum cedrina laquearia nostra cyp- 1:16 The beams of our house are cedar, and our rafters of

fir.

ressina

2:1 ego flos campi et lilium conuallium 2:2 sicut lilium 2:1 I am the rose of Sharon, and the lily of the valleys.

2:2 As the lily among thorns, so is my love among the daughters.

inter spinas sic amica mea inter filias 2:3 sicut malum in-

2:3 As the apple tree among the trees of the wood, so is my beloved among the sons. I sat down under his shadow

ter ligna siluarum sic dilectus meus inter filios sub umbra _{with great delight, and his fruit was sweet to my taste.} illius quam desideraueram sedi et fructus eius dulcis gutturi meo 2:4 introduxit me in cellam uinariam ordinauit in 2:4 He brought me to the banqueting house, and his banner

over me was love.

me caritatem 2:5 fulcite me floribus stipate me malis quia 2:5 Stay me with flagons, comfort me with apples: for I

am sick of love.

amore langueo 2:6 leua eius sub capite meo et dextera il- 2:6 His left hand is under my head, and his right hand doth

embrace me.

lius amplexabitur me 2:7 adiuro uos filiae Hierusalem per 2:7 I charge you, O ye daughters of Jerusalem, by the roes,

and by the hinds of the field, that ye stir not up, nor awake my love, till he please.

Biblia Sacra Vulgata

capreas ceruosque camporum ne suscitetis neque euigilare

2:8 The voice of my beloved! behold, he cometh leaping faciatis dilectam quoadusque ipsa uelit 2:8 uox dilecti mei

upon the mountains, skipping upon the hills.

2:9 My beloved is like a roe or a young hart: behold, he ecce iste uenit saliens in montibus transiliens colles 2:9 si-

standeth behind our wall, he looketh forth at the windows, shewing himself through the lattice.

milis est dilectus meus capreae hinuloque ceruorum en ipse stat post parietem nostrum despiciens per fenestras pros-

2:10 My beloved spake, and said unto me, Rise up, my piciens per cancellos 2:10 et dilectus meus loquitur mihi

love, my fair one, and come away. 2:11 For, lo, the winter is past, the rain is over and gone; surge propera amica mea formonsa mea et ueni 2:11 iam

2:12 The flowers appear on the earth; the time of the enim hiemps transiit imber abiit et recessit 2:12 flores appa-

singing of birds is come, and the voice of the turtle is heard in our land;

ruerunt in terra tempus putationis aduenit uox turturis audita

2:13 The fig tree putteth forth her green figs, and the vines est in terra nostra 2:13 ficus protulit grossos suos uineae flo-

with the tender grape give a good smell. Arise, my love, my fair one, and come away.

rent dederunt odorem surge amica mea speciosa mea et ueni

2:14 O my dove, that art in the clefts of the rock, in the se- 2:14 columba mea in foraminibus petrae in cauerna macecret places of the stairs, let me see thy countenance, let me hear thy voice; for sweet is thy voice, and thy countenance

is comely. riae ostende mihi faciem tuam sonet uox tua in auribus meis

2:15 Take us the foxes, the little foxes, that spoil the vines: uox enim tua dulcis et facies tua decora 2:15 capite nobis for our vines have tender grapes.

uulpes uulpes paruulas quae demoliuntur uineas nam uinea

2:16 My beloved is mine, and I am his: he feedeth among nostra floruit 2:16 dilectus meus mihi et ego illi qui pasci-

the lilies.

2:17 Until the day break, and the shadows flee away, turn, tur inter lilia 2:17 donec adspiret dies et inclinentur umbrae

my beloved, and be thou like a roe or a young hart upon the mountains of Bether.

reuertere similis esto dilecte mi capreae aut hinulo ceruorum super montes Bether

3:1 By night on my bed I sought him whom my soul 3:1 in lectulo meo per noctes quaesiui quem diligit anima

loveth: I sought him, but I found him not. 3:2 I will rise now, and go about the city in the streets, and mea quaesiui illum et non inueni 3:2 surgam et circuibo in the broad ways I will seek him whom my soul loveth: I sought him, but I found him not.

ciuitatem per uicos et plateas quaeram quem diligit anima

3:3 The watchmen that go about the city found me: to mea quaesiui illum et non inueni 3:3 inuenerunt me uigiles

whom I said, Saw ye him whom my soul loveth?

qui custodiunt ciuitatem num quem dilexit anima mea uidi-

3:4 It was but a little that I passed from them, but I found stis 3:4 paululum cum pertransissem eos inueni quem dili-

him whom my soul loveth: I held him, and would not let him go, until I had brought him into my mother’s house,

and into the chamber of her that conceived me. git anima mea tenui eum nec dimittam donec introducam il-

lum in domum matris meae et in cubiculum genetricis meae

3:5 I charge you, O ye daughters of Jerusalem, by the roes, 3:5 adiuro uos filiae Hierusalem per capreas ceruosque cam-

and by the hinds of the field, that ye stir not up, nor awake my love, till he please.

porum ne suscitetis neque euigilare faciatis dilectam donec

3:6 Who is this that cometh out of the wilderness like ipsa uelit 3:6 quae est ista quae ascendit per desertum sicut

pillars of smoke, perfumed with myrrh and frankincense, with all powders of the merchant?

uirgula fumi ex aromatibus murrae et turis et uniuersi pulue-

3:7 Behold his bed, which is Solomon’s; threescore valiant ris pigmentarii 3:7 en lectulum Salomonis sexaginta fortes

men are about it, of the valiant of Israel.

3:8 They all hold swords, being expert in war: every man ambiunt ex fortissimis Israhel 3:8 omnes tenentes gladios

hath his sword upon his thigh because of fear in the night.

Canticum Canticorum

et ad bella doctissimi uniuscuiusque ensis super femur suum propter timores nocturnos 3:9 ferculum fecit sibi rex Sa- 3:9 King Solomon made himself a chariot of the wood of

Lebanon.

lomon de lignis Libani 3:10 columnas eius fecit argenteas 3:10 He made the pillars thereof of silver, the bot-

tom thereof of gold, the covering of it of purple, the midst thereof being paved with love, for the daughters of

reclinatorium aureum ascensum purpureum media caritate Jerusalem. constrauit propter filias Hierusalem 3:11 egredimini et ui- 3:11 Go forth, O ye daughters of Zion, and behold king

Solomon with the crown wherewith his mother crowned him in the day of his espousals, and in the day of the glad-

dete filiae Sion regem Salomonem in diademate quo corona- ness of his heart. uit eum mater sua in die disponsionis illius et in die laetitiae cordis eius

4:1 quam pulchra es amica mea quam pulchra es oculi tui 4:1 Behold, thou art fair, my love; behold, thou art fair; thou hast doves’ eyes within thy locks: thy hair is as a flock of goats, that appear from mount Gilead.

columbarum absque eo quod intrinsecus latet capilli tui sicut greges caprarum quae ascenderunt de monte Galaad 4:2 4:2 Thy teeth are like a flock of sheep that are even shorn,

which came up from the washing; whereof every one bear twins, and none is barren among them.

dentes tui sicut greges tonsarum quae ascenderunt de lauacro omnes gemellis fetibus et sterilis non est inter eas 4:3 sicut 4:3 Thy lips are like a thread of scarlet, and thy speech

is comely: thy temples are like a piece of a pomegranate within thy locks.

uitta coccinea labia tua et eloquium tuum dulce sicut fragmen mali punici ita genae tuae absque eo quod intrinsecus latet 4:4 sicut turris Dauid collum tuum quae aedificata est 4:4 Thy neck is like the tower of David builded for an

armoury, whereon there hang a thousand bucklers, all shields of mighty men.

cum propugnaculis mille clypei pendent ex ea omnis armatura fortium 4:5 duo ubera tua sicut duo hinuli capreae ge- 4:5 Thy two breasts are like two young roes that are twins,

which feed among the lilies.

melli qui pascuntur in liliis 4:6 donec adspiret dies et incli- 4:6 Until the day break, and the shadows flee away, I will

get me to the mountain of myrrh, and to the hill of frankincense.

nentur umbrae uadam ad montem murrae et ad collem turis 4:7 tota pulchra es amica mea et macula non est in te 4:8 4:7 Thou art all fair, my love; there is no spot in thee.

4:8 Come with me from Lebanon, my spouse, with me from Lebanon: look from the top of Amana, from the top of Shenir and Hermon, from the lions’ dens, from the mountains of the leopards.

ueni de Libano sponsa ueni de Libano ueni coronaberis de capite Amana de uertice Sanir et Hermon de cubilibus leonum de montibus pardorum 4:9 uulnerasti cor meum soror 4:9 Thou hast ravished my heart, my sister, my spouse;

thou hast ravished my heart with one of thine eyes, with one chain of thy neck.

mea sponsa uulnerasti cor meum in uno oculorum tuorum et in uno crine colli tui 4:10 quam pulchrae sunt mammae tuae 4:10 How fair is thy love, my sister, my spouse! how much

better is thy love than wine! and the smell of thine ointments than all spices!

soror mea sponsa pulchriora ubera tua uino et odor unguentorum tuorum super omnia aromata 4:11 fauus distillans la- 4:11 Thy lips, O my spouse, drop as the honeycomb:

honey and milk are under thy tongue; and the smell of thy garments is like the smell of Lebanon.

bia tua sponsa mel et lac sub lingua tua et odor uestimentorum tuorum sicut odor turis 4:12 hortus conclusus soror 4:12 A garden inclosed is my sister, my spouse; a spring

shut up, a fountain sealed.

mea sponsa hortus conclusus fons signatus 4:13 emissiones 4:13 Thy plants are an orchard of pomegranates, with

pleasant fruits; camphire, with spikenard,

tuae paradisus malorum punicorum cum pomorum fructibus cypri cum nardo 4:14 nardus et crocus fistula et cinnamo- 4:14 Spikenard and saffron; calamus and cinnamon, with

all trees of frankincense; myrrh and aloes, with all the chief spices:

Biblia Sacra Vulgata

mum cum uniuersis lignis Libani murra et aloe cum omni-

4:15 A fountain of gardens, a well of living waters, and bus primis unguentis 4:15 fons hortorum puteus aquarum

streams from Lebanon.

4:16 Awake, O north wind; and come, thou south; blow uiuentium quae fluunt impetu de Libano 4:16 surge aquilo

upon my garden, that the spices thereof may flow out. Let my beloved come into his garden, and eat his pleas-

ant fruits. et ueni auster perfla hortum meum et fluant aromata illius

5:1 I am come into my garden, my sister, my spouse: I 5:1 ueniat dilectus meus in hortum suum et comedat fruc-

have gathered my myrrh with my spice; I have eaten my honeycomb with my honey; I have drunk my wine with my milk: eat, O friends; drink, yea, drink abundantly, O beloved.

tum pomorum suorum ueni in hortum meum soror mea sponsa messui murram meam cum aromatibus meis comedi fauum cum melle meo bibi uinum meum cum lacte meo co-

5:2 I sleep, but my heart waketh: it is the voice of my medite amici bibite et inebriamini carissimi 5:2 ego dormio

beloved that knocketh, saying, Open to me, my sister, my love, my dove, my undefiled: for my head is filled with

dew, and my locks with the drops of the night. et cor meum uigilat uox dilecti mei pulsantis aperi mihi soror mea amica mea columba mea inmaculata mea quia caput meum plenum est rore et cincinni mei guttis noctium

5:3 I have put off my coat; how shall I put it on? I have 5:3 expoliaui me tunica mea quomodo induar illa laui pe-

washed my feet; how shall I defile them?

5:4 My beloved put in his hand by the hole of the door, des meos quomodo inquinabo illos 5:4 dilectus meus misit

and my bowels were moved for him.

manum suam per foramen et uenter meus intremuit ad tac-

5:5 I rose up to open to my beloved; and my hands dropped tum eius 5:5 surrexi ut aperirem dilecto meo manus meae

with myrrh, and my fingers with sweet smelling myrrh, upon the handles of the lock.

stillauerunt murra digiti mei pleni murra probatissima 5:6

5:6 I opened to my beloved; but my beloved had withdrawn himself, and was gone: my soul failed when he spake: I sought him, but I could not find him; I called him, but he gave me no answer.

pessulum ostii aperui dilecto meo at ille declinauerat atque transierat anima mea liquefacta est ut locutus est quaesiui et

5:7 The watchmen that went about the city found me, they non inueni illum uocaui et non respondit mihi 5:7 inuene-

smote me, they wounded me; the keepers of the walls took away my veil from me.

runt me custodes qui circumeunt ciuitatem percusserunt me uulnerauerunt me tulerunt pallium meum mihi custodes mu-

5:8 I charge you, O daughters of Jerusalem, if ye find my rorum 5:8 adiuro uos filiae Hierusalem si inueneritis dilec-

beloved, that ye tell him, that I am sick of love.

5:9 What is thy beloved more than another beloved, O thou tum meum ut nuntietis ei quia amore langueo 5:9 qualis est

fairest among women? what is thy beloved more than another beloved, that thou dost so charge us?

dilectus tuus ex dilecto o pulcherrima mulierum qualis est

5:10 My beloved is white and ruddy, the chiefest among dilectus tuus ex dilecto quia sic adiurasti nos 5:10 dilectus

ten thousand. 5:11 His head is as the most fine gold, his locks are bushy, meus candidus et rubicundus electus ex milibus 5:11 caand black as a raven.

put eius aurum optimum comae eius sicut elatae palmarum

5:12 His eyes are as the eyes of doves by the rivers of nigrae quasi coruus 5:12 oculi eius sicut columbae super ri-

waters, washed with milk, and fitly set.

uulos aquarum quae lacte sunt lotae et resident iuxta fluenta

5:13 His cheeks are as a bed of spices, as sweet flowers: plenissima 5:13 genae illius sicut areolae aromatum con-

his lips like lilies, dropping sweet smelling myrrh.

sitae a pigmentariis labia eius lilia distillantia murram pri-

5:14 His hands are as gold rings set with the beryl: his mam 5:14 manus illius tornatiles aureae plenae hyacinthis

belly is as bright ivory overlaid with sapphires.

Canticum Canticorum

uenter eius eburneus distinctus sapphyris 5:15 crura illius 5:15 His legs are as pillars of marble, set upon sockets of

fine gold: his countenance is as Lebanon, excellent as the cedars.

columnae marmoreae quae fundatae sunt super bases aureas species eius ut Libani electus ut cedri 5:16 guttur illius sua- 5:16 His mouth is most sweet: yea, he is altogether lovely.

This is my beloved, and this is my friend, O daughters of Jerusalem.

uissimum et totus desiderabilis talis est dilectus meus et iste est amicus meus filiae Hierusalem 5:17 quo abiit dilectus 5:17 Whither is thy beloved gone, O thou fairest among

women? whither is thy beloved turned aside? that we may seek him with thee.

tuus o pulcherrima mulierum quo declinauit dilectus tuus et quaeremus eum tecum

6:1 dilectus meus descendit in hortum suum ad areolam 6:1 My beloved is gone down into his garden, to the beds

of spices, to feed in the gardens, and to gather lilies.

aromatis ut pascatur in hortis et lilia colligat 6:2 ego di- 6:2 I am my beloved’s, and my beloved is mine: he feedeth

among the lilies.

lecto meo et dilectus meus mihi qui pascitur inter lilia 6:3 6:3 Thou art beautiful, O my love, as Tirzah, comely as

Jerusalem, terrible as an army with banners.

pulchra es amica mea suauis et decora sicut Hierusalem terribilis ut castrorum acies ordinata 6:4 auerte oculos tuos a 6:4 Turn away thine eyes from me, for they have overcome

me: thy hair is as a flock of goats that appear from Gilead.

me quia ipsi me auolare fecerunt capilli tui sicut grex caprarum quae apparuerunt de Galaad 6:5 dentes tui sicut grex 6:5 Thy teeth are as a flock of sheep which go up from the

washing, whereof every one beareth twins, and there is not one barren among them.

ouium quae ascenderunt de lauacro omnes gemellis fetibus et sterilis non est in eis 6:6 sicut cortex mali punici genae 6:6 As a piece of a pomegranate are thy temples within thy

locks.

tuae absque occultis tuis 6:7 sexaginta sunt reginae et octo- 6:7 There are threescore queens, and fourscore concu-

bines, and virgins without number.

ginta concubinae et adulescentularum non est numerus 6:8 6:8 My dove, my undefiled is but one; she is the only one

of her mother, she is the choice one of her that bare her. The daughters saw her, and blessed her; yea, the queens

una est columba mea perfecta mea una est matris suae electa and the concubines, and they praised her. genetrici suae uiderunt illam filiae et beatissimam praedicauerunt reginae et concubinae et laudauerunt eam 6:9 quae 6:9 Who is she that looketh forth as the morning, fair as

the moon, clear as the sun, and terrible as an army with banners?

est ista quae progreditur quasi aurora consurgens pulchra ut luna electa ut sol terribilis ut acies ordinata 6:10 descendi 6:10 I went down into the garden of nuts to see the fruits

of the valley, and to see whether the vine flourished and the pomegranates budded.

ad hortum nucum ut uiderem poma conuallis ut inspicerem si floruisset uinea et germinassent mala punica 6:11 nes- 6:11 Or ever I was aware, my soul made me like the char-

iots of Amminadib.

ciui anima mea conturbauit me propter quadrigas Aminadab 6:12 reuertere reuertere Sulamitis reuertere reuertere ut in- 6:12 Return, return, O Shulamite; return, return, that we

may look upon thee.

tueamur te

7:1 quid uidebis in Sulamiten nisi choros castrorum quam 7:1 What will ye see in the Shulamite? As it were the

company of two armies. How beautiful are thy feet with shoes, O prince’s daughter! the joints of thy thighs are like

pulchri sunt gressus tui in calciamentis filia principis iunc- jewels, the work of the hands of a cunning workman. tura feminum tuorum sicut monilia quae fabricata sunt manu artificis 7:2 umbilicus tuus crater tornatilis numquam indi- 7:2 Thy navel is like a round goblet, which wanteth not

liquor: thy belly is like an heap of wheat set about with lilies.

gens poculis uenter tuus sicut aceruus tritici uallatus liliis

Biblia Sacra Vulgata

7:3 Thy two breasts are like two young roes that are twins. 7:3 duo ubera tua sicut duo hinuli gemelli capreae 7:4 col-

7:4 Thy neck is as a tower of ivory; thine eyes like the fishpools in Heshbon, by the gate of Bathrabbim: thy nose is as the tower of Lebanon which looketh toward Damascus.

lum tuum sicut turris eburnea oculi tui sicut piscinae in Esebon quae sunt in porta filiae multitudinis nasus tuus sicut tur-

7:5 Thine head upon thee is like Carmel, and the hair of ris Libani quae respicit contra Damascum 7:5 caput tuum

thine head like purple; the king is held in the galleries.

ut Carmelus et comae capitis tui sicut purpura regis uincta

7:6 How fair and how pleasant art thou, O love, for de- canalibus 7:6 quam pulchra es et quam decora carissima in

lights!

7:7 This thy stature is like to a palm tree, and thy breasts deliciis 7:7 statura tua adsimilata est palmae et ubera tua

to clusters of grapes.

7:8 I said, I will go up to the palm tree, I will take hold botris 7:8 dixi ascendam in palmam adprehendam fructus

of the boughs thereof: now also thy breasts shall be as clusters of the vine, and the smell of thy nose like apples;

eius et erunt ubera tua sicut botri uineae et odor oris tui sicut

7:9 And the roof of thy mouth like the best wine for my malorum 7:9 guttur tuum sicut uinum optimum dignum di-

beloved, that goeth down sweetly, causing the lips of those that are asleep to speak.

lecto meo ad potandum labiisque et dentibus illius ruminan-

7:10 I am my beloved’s, and his desire is toward me. dum 7:10 ego dilecto meo et ad me conuersio eius 7:11

7:11 Come, my beloved, let us go forth into the field; let us lodge in the villages.

ueni dilecte mi egrediamur in agrum commoremur in uillis

7:12 Let us get up early to the vineyards; let us see if 7:12 mane surgamus ad uineas uideamus si floruit uinea si

the vine flourish, whether the tender grape appear, and the pomegranates bud forth: there will I give thee my loves.

flores fructus parturiunt si floruerunt mala punica ibi dabo

7:13 The mandrakes give a smell, and at our gates are all tibi ubera mea 7:13 mandragorae dederunt odorem in por-

manner of pleasant fruits, new and old, which I have laid up for thee, O my beloved.

tis nostris omnia poma noua et uetera dilecte mi seruaui tibi

8:1 O that thou wert as my brother, that sucked the breasts 8:1 quis mihi det te fratrem meum sugentem ubera matris of my mother! when I should find thee without, I would kiss thee; yea, I should not be despised.

meae ut inueniam te foris et deosculer et iam me nemo despi-

8:2 I would lead thee, and bring thee into my mother’s ciat 8:2 adprehendam te et ducam in domum matris meae house, who would instruct me: I would cause thee to drink of spiced wine of the juice of my pomegranate.

ibi me docebis et dabo tibi poculum ex uino condito et mus-

8:3 His left hand should be under my head, and his right tum malorum granatorum meorum 8:3 leua eius sub capite

hand should embrace me. 8:4 I charge you, O daughters of Jerusalem, that ye stir not meo et dextera illius amplexabitur me 8:4 adiuro uos filiae up, nor awake my love, until he please.

Hierusalem ne suscitetis et euigilare faciatis dilectam donec

8:5 Who is this that cometh up from the wilderness, lean- ipsa uelit 8:5 quae est ista quae ascendit de deserto deliciis

ing upon her beloved? I raised thee up under the apple tree: there thy mother brought thee forth: there she

brought thee forth that bare thee. affluens et nixa super dilectum suum sub arbore malo susci-

taui te ibi corrupta est mater tua ibi uiolata est genetrix tua

8:6 Set me as a seal upon thine heart, as a seal upon thine 8:6 pone me ut signaculum super cor tuum ut signaculum su-

arm: for love is strong as death; jealousy is cruel as the grave: the coals thereof are coals of fire, which hath a most

vehement flame. per brachium tuum quia fortis est ut mors dilectio dura sicut

inferus aemulatio lampades eius lampades ignis atque flam-

8:7 Many waters cannot quench love, neither can the marum 8:7 aquae multae non poterunt extinguere caritatem

floods drown it: if a man would give all the substance of his house for love, it would utterly be contemned.

nec flumina obruent illam si dederit homo omnem substantiam domus suae pro dilectione quasi nihil despicient eum

Canticum Canticorum

8:8 soror nostra parua et ubera non habet quid faciemus so- 8:8 We have a little sister, and she hath no breasts: what

shall we do for our sister in the day when she shall be spoken for?

rori nostrae in die quando adloquenda est 8:9 si murus est

8:9 If she be a wall, we will build upon her a palace of silver: and if she be a door, we will inclose her with boards of cedar.

aedificemus super eum propugnacula argentea si ostium est conpingamus illud tabulis cedrinis 8:10 ego murus et ubera 8:10 I am a wall, and my breasts like towers: then was I in

his eyes as one that found favour.

mea sicut turris ex quo facta sum coram eo quasi pacem repperiens 8:11 uinea fuit Pacifico in ea quae habet populos 8:11 Solomon had a vineyard at Baalhamon; he let out the

vineyard unto keepers; every one for the fruit thereof was to bring a thousand pieces of silver.

tradidit eam custodibus uir adfert pro fructu eius mille argenteos 8:12 uinea mea coram me est mille tui Pacifice et 8:12 My vineyard, which is mine, is before me: thou, O

Solomon, must have a thousand, and those that keep the fruit thereof two hundred.

ducenti his qui custodiunt fructus eius 8:13 quae habitas in

8:13 Thou that dwellest in the gardens, the companions hearken to thy voice: cause me to hear it.

hortis amici auscultant fac me audire uocem tuam 8:14 fuge 8:14 Make haste, my beloved, and be thou like to a roe or

to a young hart upon the mountains of spices.

dilecte mi et adsimilare capreae hinuloque ceruorum super montes aromatum

Biblia Sacra Vulgata

CMLXII

Prophetia Isaiae

1:1 uisio Isaiae filii Amos quam uidit super Iudam et Hie- 1:1 The vision of Isaiah the son of Amoz, which he saw

concerning Judah and Jerusalem in the days of Uzziah, Jotham, Ahaz, and Hezekiah, kings of Judah.

rusalem in diebus Oziae Ioatham Ahaz Ezechiae regum Iuda 1:2 audite caeli et auribus percipe terra quoniam Dominus 1:2 Hear, O heavens, and give ear, O earth: for the LORD

hath spoken, I have nourished and brought up children, and they have rebelled against me.

locutus est filios enutriui et exaltaui ipsi autem spreuerunt me 1:3 cognouit bos possessorem suum et asinus praesepe 1:3 The ox knoweth his owner, and the ass his master’s

crib: but Israel doth not know, my people doth not consider.

domini sui Israhel non cognouit populus meus non intellexit 1:4 uae genti peccatrici populo graui iniquitate semini ne- 1:4 Ah sinful nation, a people laden with iniquity, a seed of

evildoers, children that are corrupters: they have forsaken the LORD, they have provoked the Holy One of Israel unto

quam filiis sceleratis dereliquerunt Dominum blasphemaue- anger, they are gone away backward. runt Sanctum Israhel abalienati sunt retrorsum 1:5 super 1:5 Why should ye be stricken any more? ye will revolt

more and more: the whole head is sick, and the whole heart faint.

quo percutiam uos ultra addentes praeuaricationem omne caput languidum et omne cor maerens 1:6 a planta pedis 1:6 From the sole of the foot even unto the head there is

no soundness in it; but wounds, and bruises, and putrifying sores: they have not been closed, neither bound up, neither

usque ad uerticem non est in eo sanitas uulnus et liuor et mollified with ointment. plaga tumens non est circumligata nec curata medicamine neque fota oleo 1:7 terra uestra deserta ciuitates uestrae 1:7 Your country is desolate, your cities are burned with

fire: your land, strangers devour it in your presence, and it is desolate, as overthrown by strangers.

succensae igni regionem uestram coram uobis alieni deuorant et desolabitur sicut in uastitate hostili 1:8 et derelin- 1:8 And the daughter of Zion is left as a cottage in a vine-

yard, as a lodge in a garden of cucumbers, as a besieged city.

quetur filia Sion ut umbraculum in uinea et sicut tugurium in cucumerario sicut ciuitas quae uastatur 1:9 nisi Domi- 1:9 Except the LORD of hosts had left unto us a very small

remnant, we should have been as Sodom, and we should have been like unto Gomorrah.

nus exercituum reliquisset nobis semen quasi Sodoma fuissemus et quasi Gomorra similes essemus 1:10 audite uer- 1:10 Hear the word of the LORD, ye rulers of Sodom; give

ear unto the law of our God, ye people of Gomorrah.

bum Domini principes Sodomorum percipite auribus legem Dei nostri populus Gomorrae 1:11 quo mihi multitudinem 1:11 To what purpose is the multitude of your sacrifices

unto me? saith the LORD: I am full of the burnt offerings of rams, and the fat of fed beasts; and I delight not in the

uictimarum uestrarum dicit Dominus plenus sum holocausta blood of bullocks, or of lambs, or of he goats. arietum et adipem pinguium et sanguinem uitulorum et agnorum et hircorum nolui 1:12 cum ueneritis ante conspec- 1:12 When ye come to appear before me, who hath re-

quired this at your hand, to tread my courts?

tum meum quis quaesiuit haec de manibus uestris ut ambularetis in atriis meis 1:13 ne adferatis ultra sacrificium frus- 1:13 Bring no more vain oblations; incense is an abomi-

nation unto me; the new moons and sabbaths, the calling of assemblies, I cannot away with; it is iniquity, even the

tra incensum abominatio est mihi neomeniam et sabbatum solemn meeting. et festiuitates alias non feram iniqui sunt coetus uestri 1:14 1:14 Your new moons and your appointed feasts my soul

hateth: they are a trouble unto me; I am weary to bear them.

kalendas uestras et sollemnitates uestras odiuit anima mea facta sunt mihi molesta laboraui sustinens 1:15 et cum ex- 1:15 And when ye spread forth your hands, I will hide

mine eyes from you: yea, when ye make many prayers, I will not hear: your hands are full of blood.

tenderitis manus uestras auertam oculos meos a uobis et cum multiplicaueritis orationem non audiam manus uestrae san-

Biblia Sacra Vulgata

1:16 Wash you, make you clean; put away the evil of your guine plenae sunt 1:16 lauamini mundi estote auferte ma-

doings from before mine eyes; cease to do evil;

lum cogitationum uestrarum ab oculis meis quiescite agere

1:17 Learn to do well; seek judgment, relieve the op- peruerse 1:17 discite benefacere quaerite iudicium subue-

pressed, judge the fatherless, plead for the widow.

1:18 Come now, and let us reason together, saith the nite oppresso iudicate pupillo defendite uiduam 1:18 et ue-

LORD: though your sins be as scarlet, they shall be as white as snow; though they be red like crimson, they shall

be as wool. nite et arguite me dicit Dominus si fuerint peccata uestra

ut coccinum quasi nix dealbabuntur et si fuerint rubra quasi

1:19 If ye be willing and obedient, ye shall eat the good of uermiculus uelut lana erunt 1:19 si uolueritis et audieritis

the land:

1:20 But if ye refuse and rebel, ye shall be devoured with bona terrae comedetis 1:20 quod si nolueritis et me prouo-

the sword: for the mouth of the LORD hath spoken it.

caueritis ad iracundiam gladius deuorabit uos quia os Do-

1:21 How is the faithful city become an harlot! it was full mini locutum est 1:21 quomodo facta est meretrix ciuitas of judgment; righteousness lodged in it; but now murderers.

fidelis plena iudicii iustitia habitauit in ea nunc autem ho-

1:22 Thy silver is become dross, thy wine mixed with wa- micidae 1:22 argentum tuum uersum est in scoriam uinum ter:

1:23 Thy princes are rebellious, and companions of tuum mixtum est aqua 1:23 principes tui infideles socii fu-

thieves: every one loveth gifts, and followeth after rewards: they judge not the fatherless, neither doth the cause

of the widow come unto them. rum omnes diligunt munera sequuntur retributiones pupillo 1:24 Therefore saith the LORD, the LORD of hosts, the non iudicant et causa uiduae non ingreditur ad eos 1:24 pro-

mighty One of Israel, Ah, I will ease me of mine adversaries, and avenge me of mine enemies:

pter hoc ait Dominus exercituum Fortis Israhel heu consola-

1:25 And I will turn my hand upon thee, and purely purge bor super hostibus meis et uindicabor de inimicis meis 1:25

away thy dross, and take away all thy tin:

et conuertam manum meam ad te et excoquam ad purum

1:26 And I will restore thy judges as at the first, and thy scoriam tuam et auferam omne stagnum tuum 1:26 et rescounsellors as at the beginning: afterward thou shalt be called, The city of righteousness, the faithful city.

tituam iudices tuos ut fuerunt prius et consiliarios tuos sicut

1:27 Zion shall be redeemed with judgment, and her con- antiquitus post haec uocaberis ciuitas iusti urbs fidelis 1:27

verts with righteousness.

1:28 And the destruction of the transgressors and of the Sion in iudicio rediietur et reducent eam in iustitia 1:28 et

sinners shall be together, and they that forsake the LORD shall be consumed.

conteret scelestos et peccatores simul et qui dereliquerunt

1:29 For they shall be ashamed of the oaks which ye have Dominum consumentur 1:29 confundentur enim ab idolis

desired, and ye shall be confounded for the gardens that ye have chosen.

quibus sacrificauerunt et erubescetis super hortis quos ele-

1:30 For ye shall be as an oak whose leaf fadeth, and as a geratis 1:30 cum fueritis uelut quercus defluentibus foliis

garden that hath no water.

1:31 And the strong shall be as tow, and the maker of it as et uelut hortus absque aqua 1:31 et erit fortitudo uestra ut

a spark, and they shall both burn together, and none shall quench them.

fauilla stuppae et opus uestrum quasi scintilla et succendetur utrumque simul et non erit qui extinguat

2:1 The word that Isaiah the son of Amoz saw concerning 2:1 uerbum quod uidit Isaias filius Amos super Iudam Judah and Jerusalem.

2:2 And it shall come to pass in the last days, that the et Hierusalem 2:2 et erit in nouissimis diebus praeparatus

mountain of the LORD’s house shall be established in the top of the mountains, and shall be exalted above the hills;

and all nations shall flow unto it. mons domus Domini in uertice montium et eleuabitur super 2:3 And many people shall go and say, Come ye, and let colles et fluent ad eum omnes gentes 2:3 et ibunt populi

us go up to the mountain of the LORD, to the house of the God of Jacob; and he will teach us of his ways, and we will walk in his paths: for out of Zion shall go forth the law, and the word of the LORD from Jerusalem.

Prophetia Isaiae

multi et dicent uenite et ascendamus ad montem Domini et ad domum Dei Iacob et docebit nos uias suas et ambulabimus in semitis eius quia de Sion exibit lex et uerbum Do-

mini de Hierusalem 2:4 et iudicabit gentes et arguet po- 2:4 And he shall judge among the nations, and shall rebuke many people: and they shall beat their swords into plowshares, and their spears into pruninghooks: nation shall not lift up sword against nation, neither shall they learn war any more.

pulos multos et conflabunt gladios suos in uomeres et lanceas suas in falces non leuabit gens contra gentem gladium nec exercebuntur ultra ad proelium 2:5 domus Iacob uenite 2:5 O house of Jacob, come ye, and let us walk in the light

of the LORD.

et ambulemus in lumine Domini 2:6 proiecisti enim popu- 2:6 Therefore thou hast forsaken thy people the house

of Jacob, because they be replenished from the east, and are soothsayers like the Philistines, and they please them-

lum tuum domum Iacob quia repleti sunt ut olim et augu- selves in the children of strangers. res habuerunt ut Philisthim et pueris alienis adheserunt 2:7 2:7 Their land also is full of silver and gold, neither is there

any end of their treasures; their land is also full of horses, neither is there any end of their chariots:

repleta est terra argento et auro et non est finis thesaurorum eius 2:8 et repleta est terra eius equis et innumerabi- 2:8 Their land also is full of idols; they worship the work

of their own hands, that which their own fingers have made:

les quadrigae eius et repleta est terra eius idolis opus manuum suarum adorauerunt quod fecerunt digiti eorum 2:9 2:9 And the mean man boweth down, and the great man

humbleth himself: therefore forgive them not.

et incuruauit se homo et humiliatus est uir ne ergo dimit-

tas eis 2:10 ingredere in petram abscondere fossa humo a 2:10 Enter into the rock, and hide thee in the dust, for fear

of the LORD, and for the glory of his majesty.

facie timoris Domini et a gloria maiestatis eius 2:11 oculi 2:11 The lofty looks of man shall be humbled, and the

haughtiness of men shall be bowed down, and the LORD alone shall be exalted in that day.

sublimis hominis humiliati sunt et incuruabitur altitudo uirorum exaltabitur autem Dominus solus in die illa 2:12 quia 2:12 For the day of the LORD of hosts shall be upon every

one that is proud and lofty, and upon every one that is lifted up; and he shall be brought low:

dies Domini exercituum super omnem superbum et excelsum et super omnem arrogantem et humiliabitur 2:13 et su- 2:13 And upon all the cedars of Lebanon, that are high and

lifted up, and upon all the oaks of Bashan,

per omnes cedros Libani sublimes et erectas et super omnes quercus Basan 2:14 et super omnes montes excelsos et 2:14 And upon all the high mountains, and upon all the

hills that are lifted up,

super omnes colles eleuatos 2:15 et super omnem turrem 2:15 And upon every high tower, and upon every fenced

wall,

excelsam et super omnem murum munitum 2:16 et super 2:16 And upon all the ships of Tarshish, and upon all

pleasant pictures.

omnes naues Tharsis et super omne quod uisu pulchrum est 2:17 et incuruabitur sublimitas hominum et humiliabitur alti- 2:17 And the loftiness of man shall be bowed down, and

the haughtiness of men shall be made low: and the LORD alone shall be exalted in that day.

tudo uirorum et eleuabitur Dominus solus in die illa 2:18 et

2:18 And the idols he shall utterly abolish.

idola penitus conterentur 2:19 et introibunt in speluncas pe- 2:19 And they shall go into the holes of the rocks, and into

the caves of the earth, for fear of the LORD, and for the glory of his majesty, when he ariseth to shake terribly the

trarum et in uoragines terrae a facie formidinis Domini et a earth. gloria maiestatis eius cum surrexerit percutere terram 2:20 2:20 In that day a man shall cast his idols of silver, and

his idols of gold, which they made each one for himself to worship, to the moles and to the bats;

in die illa proiciet homo idola argenti sui et simulacra auri sui quae fecerat sibi ut adoraret talpas et uespertiliones 2:21 2:21 To go into the clefts of the rocks, and into the tops of

the ragged rocks, for fear of the LORD, and for the glory of his majesty, when he ariseth to shake terribly the earth.

Biblia Sacra Vulgata

et ingredietur fissuras petrarum et cauernas saxorum a facie formidinis Domini et a gloria maiestatis eius cum surrexerit

2:22 Cease ye from man, whose breath is in his nostrils: percutere terram 2:22 quiescite ergo ab homine cuius spiri-

for wherein is he to be accounted of ?

tus in naribus eius quia excelsus reputatus est ipse

3:1 For, behold, the Lord, the LORD of hosts, doth take 3:1 ecce enim Dominator Deus exercituum auferet ab Hie-

away from Jerusalem and from Judah the stay and the staff, the whole stay of bread, and the whole stay of water.

rusalem et ab Iuda ualidum et fortem omne robur panis et

3:2 The mighty man, and the man of war, the judge, and omne robur aquae 3:2 fortem et uirum bellatorem iudi-

the prophet, and the prudent, and the ancient, 3:3 The captain of fifty, and the honourable man, and the cem et prophetam et ariolum et senem 3:3 principem sucounsellor, and the cunning artificer, and the eloquent orator.

per quinquaginta et honorabilem uultu et consiliarium sa-

3:4 And I will give children to be their princes, and babes pientem de architectis et prudentem eloquii mystici 3:4 et

shall rule over them.

dabo pueros principes eorum et effeminati dominabuntur

3:5 And the people shall be oppressed, every one by an- eis 3:5 et inruet populus uir ad uirum unusquisque ad pro-

other, and every one by his neighbour: the child shall behave himself proudly against the ancient, and the base

against the honourable. ximum suum tumultuabitur puer contra senem et ignobilis 3:6 When a man shall take hold of his brother of the house contra nobilem 3:6 adprehendet enim uir fratrem suum do-

of his father, saying, Thou hast clothing, be thou our ruler, and let this ruin be under thy hand:

mesticum patris sui uestimentum tibi est princeps esto nos-

3:7 In that day shall he swear, saying, I will not be an ter ruina autem haec sub manu tua 3:7 respondebit in die

healer; for in my house is neither bread nor clothing: make me not a ruler of the people.

illa dicens non sum medicus et in domo mea non est panis neque uestimentum nolite constituere me principem po-

3:8 For Jerusalem is ruined, and Judah is fallen: because puli 3:8 ruit enim Hierusalem et Iudas concidit quia lingua

their tongue and their doings are against the LORD, to provoke the eyes of his glory.

eorum et adinuentiones eorum contra Dominum ut prouo-

3:9 The shew of their countenance doth witness against carent oculos maiestatis eius 3:9 agnitio uultus eorum res-

them; and they declare their sin as Sodom, they hide it not. Woe unto their soul! for they have rewarded evil unto

themselves. pondit eis et peccatum suum quasi Sodomae praedicauerunt

nec absconderunt uae animae eorum quoniam reddita sunt

3:10 Say ye to the righteous, that it shall be well with him: eis mala 3:10 dicite iusto quoniam bene quoniam fructum for they shall eat the fruit of their doings. 3:11 Woe unto the wicked! it shall be ill with him: for the adinuentionum suarum comedet 3:11 uae impio in malum reward of his hands shall be given him. 3:12 As for my people, children are their oppressors, and retributio enim manuum eius fiet ei 3:12 populum meum women rule over them. O my people, they which lead thee cause thee to err, and destroy the way of thy paths.

exactores sui spoliauerunt et mulieres dominatae sunt eius popule meus qui beatum te dicunt ipsi te decipiunt et uiam

3:13 The LORD standeth up to plead, and standeth to gressuum tuorum dissipant 3:13 stat ad iudicandum Domi-

judge the people. 3:14 The LORD will enter into judgment with the ancients nus et stat ad iudicandos populos 3:14 Dominus ad iudiof his people, and the princes thereof: for ye have eaten up the vineyard; the spoil of the poor is in your houses.

cium ueniet cum senibus populi sui et principibus eius uos enim depasti estis uineam meam et rapina pauperis in domo

3:15 What mean ye that ye beat my people to pieces, and uestra 3:15 quare adteritis populum meum et facies pau-

grind the faces of the poor? saith the Lord GOD of hosts.

Prophetia Isaiae

perum commolitis dicit Dominus Deus exercituum 3:16 et 3:16 Moreover the LORD saith, Because the daughters of

Zion are haughty, and walk with stretched forth necks and wanton eyes, walking and mincing as they go, and making

dixit Dominus pro eo quod eleuatae sunt filiae Sion et am- a tinkling with their feet: bulauerunt extento collo et nutibus oculorum ibant et plaudebant ambulabant et in pedibus suis conposito gradu incedebant 3:17 decaluabit Dominus uerticem filiarum Sion et 3:17 Therefore the LORD will smite with a scab the crown

of the head of the daughters of Zion, and the LORD will discover their secret parts.

Dominus crinem earum nudabit 3:18 in die illa auferet Do-

3:18 In that day the Lord will take away the bravery of their tinkling ornaments about their feet, and their cauls, and their round tires like the moon,

minus ornatum calciamentorum et lunulas 3:19 et torques

3:19 The chains, and the bracelets, and the mufflers,

et monilia et armillas et mitras 3:20 discriminalia et peris- 3:20 The bonnets, and the ornaments of the legs, and the

headbands, and the tablets, and the earrings,

celidas et murenulas et olfactoriola et inaures 3:21 et anulos 3:21 The rings, and nose jewels, et gemmas in fronte pendentes 3:22 et mutatoria et pallia et 3:22 The changeable suits of apparel, and the mantles, and

the wimples, and the crisping pins,

linteamina et acus 3:23 et specula et sindones et uittas et 3:23 The glasses, and the fine linen, and the hoods, and

the vails.

theristra 3:24 et erit pro suaui odore fetor et pro zona funi- 3:24 And it shall come to pass, that instead of sweet smell

there shall be stink; and instead of a girdle a rent; and instead of well set hair baldness; and instead of a stomacher

culus et pro crispanti crine caluitium et pro fascia pectorali a girding of sackcloth; and burning instead of beauty. cilicium 3:25 pulcherrimi quoque uiri tui gladio cadent et 3:25 Thy men shall fall by the sword, and thy mighty in

the war.

fortes tui in proelio 3:26 et maerebunt atque lugebunt por- 3:26 And her gates shall lament and mourn; and she being

desolate shall sit upon the ground.

tae eius et desolata in terra sedebit

4:1 et adprehendent septem mulieres uirum unum in die 4:1 And in that day seven women shall take hold of one

man, saying, We will eat our own bread, and wear our own apparel: only let us be called by thy name, to take

illa dicentes panem nostrum comedemus et uestimentis no- away our reproach. stris operiemur tantummodo uocetur nomen tuum super nos aufer obprobrium nostrum 4:2 in die illa erit germen Do- 4:2 In that day shall the branch of the LORD be beautiful

and glorious, and the fruit of the earth shall be excellent and comely for them that are escaped of Israel.

mini in magnificentia et in gloria et fructus terrae sublimis et exultatio his qui saluati fuerint de Israhel 4:3 et erit omnis 4:3 And it shall come to pass, that he that is left in

Zion, and he that remaineth in Jerusalem, shall be called holy, even every one that is written among the living in

qui relictus fuerit in Sion et residuus in Hierusalem sanctus Jerusalem: uocabitur omnis qui scriptus est in uita in Hierusalem 4:4 4:4 When the Lord shall have washed away the filth of

the daughters of Zion, and shall have purged the blood of Jerusalem from the midst thereof by the spirit of judgment,

si abluerit Dominus sordem filiarum Sion et sanguinem Hie- and by the spirit of burning. rusalem lauerit de medio eius spiritu iudicii et spiritu ardoris 4:5 et creabit Dominus super omnem locum montis Sion et 4:5 And the LORD will create upon every dwelling place

of mount Zion, and upon her assemblies, a cloud and smoke by day, and the shining of a flaming fire by night:

ubi inuocatus est nubem per diem et fumum et splendorem for upon all the glory shall be a defence. ignis flammantis in nocte super omnem enim gloriam protectio 4:6 et tabernaculum erit in umbraculum diei ab aestu 4:6 And there shall be a tabernacle for a shadow in the day

time from the heat, and for a place of refuge, and for a covert from storm and from rain.

et in securitatem et absconsionem a turbine et a pluuia

5:1 cantabo dilecto meo canticum patruelis mei uineae 5:1 Now will I sing to my wellbeloved a song of my

beloved touching his vineyard. My wellbeloved hath a vineyard in a very fruitful hill:

suae uinea facta est dilecto meo in cornu filio olei 5:2 et

5:2 And he fenced it, and gathered out the stones thereof, and planted it with the choicest vine, and built a tower in the midst of it, and also made a winepress therein: and he looked that it should bring forth grapes, and it brought forth wild grapes.

CMLXVII

Biblia Sacra Vulgata

sepiuit eam et lapides elegit ex illa et plantauit eam electam et aedificauit turrem in medio eius et torcular extruxit in ea et

5:3 And now, O inhabitants of Jerusalem, and men of Ju- expectauit ut faceret uuas et fecit labruscas 5:3 nunc ergo

dah, judge, I pray you, betwixt me and my vineyard.

habitator Hierusalem et uir Iuda iudicate inter me et inter

5:4 What could have been done more to my vineyard, that uineam meam 5:4 quid est quod debui ultra facere uineae

I have not done in it? wherefore, when I looked that it should bring forth grapes, brought it forth wild grapes?

meae et non feci ei an quod expectaui ut faceret uuas et fecit

5:5 And now go to; I will tell you what I will do to my labruscas 5:5 et nunc ostendam uobis quid ego faciam ui-

vineyard: I will take away the hedge thereof, and it shall be eaten up; and break down the wall thereof, and it shall

be trodden down: neae meae auferam sepem eius et erit in direptionem diruam 5:6 And I will lay it waste: it shall not be pruned, nor maceriam eius et erit in conculcationem 5:6 et ponam eam

digged; but there shall come up briers and thorns: I will also command the clouds that they rain no rain upon it.

desertam non putabitur et non fodietur et ascendent uepres et spinae et nubibus mandabo ne pluant super eam imbrem

5:7 For the vineyard of the LORD of hosts is the house 5:7 uinea enim Domini exercituum domus Israhel et uir Iuda

of Israel, and the men of Judah his pleasant plant: and he looked for judgment, but behold oppression; for righteous-

ness, but behold a cry. germen delectabile eius et expectaui ut faceret iudicium et 5:8 Woe unto them that join house to house, that lay field ecce iniquitas et iustitiam et ecce clamor 5:8 uae qui con-

to field, till there be no place, that they may be placed alone in the midst of the earth!

iungitis domum ad domum et agrum agro copulatis usque ad terminum loci numquid habitabitis soli uos in medio ter-

5:9 In mine ears said the LORD of hosts, Of a truth many rae 5:9 in auribus meis sunt haec Domini exercituum nisi

houses shall be desolate, even great and fair, without inhabitant.

domus multae desertae fuerint grandes et pulchrae absque

5:10 Yea, ten acres of vineyard shall yield one bath, and habitatore 5:10 decem enim iuga uinearum facient lagun-

the seed of an homer shall yield an ephah.

culam unam et triginta modii sementis facient modios tres

5:11 Woe unto them that rise up early in the morning, that 5:11 uae qui consurgitis mane ad ebrietatem sectandam et

they may follow strong drink; that continue until night, till wine inflame them!

potandum usque ad uesperam ut uino aestuetis 5:12 cithara

5:12 And the harp, and the viol, the tabret, and pipe, and wine, are in their feasts: but they regard not the work of the LORD, neither consider the operation of his hands.

et lyra et tympanum et tibia et uinum in conuiuiis uestris et opus Domini non respicitis nec opera manuum eius con-

5:13 Therefore my people are gone into captivity, because sideratis 5:13 propterea captiuus ductus est populus meus

they have no knowledge: and their honourable men are famished, and their multitude dried up with thirst.

quia non habuit scientiam et nobiles eius interierunt fame et

5:14 Therefore hell hath enlarged herself, and opened her multitudo eius siti exaruit 5:14 propterea dilatauit infernus

mouth without measure: and their glory, and their multitude, and their pomp, and he that rejoiceth, shall descend

into it. animam suam et aperuit os suum absque ullo termino et de-

scendent fortes eius et populus eius et sublimes gloriosique

5:15 And the mean man shall be brought down, and the eius ad eum 5:15 et incuruabitur homo et humiliabitur uir et

mighty man shall be humbled, and the eyes of the lofty shall be humbled:

oculi sublimium deprimentur 5:16 et exaltabitur Dominus

5:16 But the LORD of hosts shall be exalted in judgment, and God that is holy shall be sanctified in righteousness.

exercituum in iudicio et Deus sanctus sanctificabitur in ius-

5:17 Then shall the lambs feed after their manner, and the titia 5:17 et pascentur agni iuxta ordinem suum et deserta

waste places of the fat ones shall strangers eat.

Prophetia Isaiae

in ubertatem uersa aduenae comedent 5:18 uae qui trahitis 5:18 Woe unto them that draw iniquity with cords of van-

ity, and sin as it were with a cart rope:

iniquitatem in funiculis uanitatis et quasi uinculum plaustri peccatum 5:19 qui dicitis festinet et cito ueniat opus eius ut 5:19 That say, Let him make speed, and hasten his work,

that we may see it: and let the counsel of the Holy One of Israel draw nigh and come, that we may know it!

uideamus et adpropiet et ueniat consilium Sancti Israhel et sciemus illud 5:20 uae qui dicitis malum bonum et bonum 5:20 Woe unto them that call evil good, and good evil;

that put darkness for light, and light for darkness; that put bitter for sweet, and sweet for bitter!

malum ponentes tenebras lucem et lucem tenebras ponentes amarum in dulce et dulce in amarum 5:21 uae qui sapien- 5:21 Woe unto them that are wise in their own eyes, and

prudent in their own sight!

tes estis in oculis uestris et coram uobismet ipsis prudentes 5:22 uae qui potentes estis ad bibendum uinum et uiri fortes 5:22 Woe unto them that are mighty to drink wine, and

men of strength to mingle strong drink:

ad miscendam ebrietatem 5:23 qui iustificatis impium pro 5:23 Which justify the wicked for reward, and take away

the righteousness of the righteous from him!

muneribus et iustitiam iusti aufertis ab eo 5:24 propter hoc 5:24 Therefore as the fire devoureth the stubble, and the

flame consumeth the chaff, so their root shall be as rottenness, and their blossom shall go up as dust: because they have cast away the law of the LORD of hosts, and despised the word of the Holy One of Israel.

sicut deuorat stipulam lingua ignis et calor flammae exurit sic radix eorum quasi fauilla erit et germen eorum ut puluis ascendet abiecerunt enim legem Domini exercituum et eloquium Sancti Israhel blasphemauerunt 5:25 ideo iratus est 5:25 Therefore is the anger of the LORD kindled against

his people, and he hath stretched forth his hand against them, and hath smitten them: and the hills did tremble, and their carcases were torn in the midst of the streets. For all this his anger is not turned away, but his hand is stretched out still.

furor Domini in populo suo et extendit manum suam super eum et percussit eum et conturbati sunt montes et facta sunt morticina eorum quasi stercus in medio platearum in omnibus his non est auersus furor eius sed adhuc manus eius extenta 5:26 et leuabit signum nationibus procul et sibilabit ad 5:26 And he will lift up an ensign to the nations from far,

and will hiss unto them from the end of the earth: and, behold, they shall come with speed swiftly:

eum de finibus terrae et ecce festinus uelociter ueniet 5:27

5:27 None shall be weary nor stumble among them; none shall slumber nor sleep; neither shall the girdle of their loins be loosed, nor the latchet of their shoes be broken:

non est deficiens neque laborans in eo non dormitabit neque dormiet neque soluetur cingulum renum eius nec rumpetur corrigia calciamenti eius 5:28 sagittae eius acutae et omnes 5:28 Whose arrows are sharp, and all their bows bent, their

horses’ hoofs shall be counted like flint, and their wheels like a whirlwind:

arcus eius extenti ungulae equorum eius ut silex et rotae eius quasi impetus tempestatis 5:29 rugitus eius ut leonis rugiet 5:29 Their roaring shall be like a lion, they shall roar like

young lions: yea, they shall roar, and lay hold of the prey, and shall carry it away safe, and none shall deliver it.

ut catuli leonum et frendet et tenebit praedam et amplexabitur et non erit qui eruat 5:30 et sonabit super eum in die 5:30 And in that day they shall roar against them like the

roaring of the sea: and if one look unto the land, behold darkness and sorrow, and the light is darkened in the heav-

illa sicut sonitus maris aspiciemus in terram et ecce tenebrae ens thereof. tribulationis et lux obtenebrata est in caligine eius

6:1 in anno quo mortuus est rex Ozias uidi Dominum se- 6:1 In the year that king Uzziah died I saw also the LORD sitting upon a throne, high and lifted up, and his train filled the temple.

dentem super solium excelsum et eleuatum et ea quae sub eo erant implebant templum 6:2 seraphin stabant super il- 6:2 Above it stood the seraphims: each one had six wings;

with twain he covered his face, and with twain he covered his feet, and with twain he did fly.

Biblia Sacra Vulgata

lud sex alae uni et sex alae alteri duabus uelabant faciem eius

6:3 And one cried unto another, and said, Holy, holy, holy, et duabus uelabant pedes eius et duabus uolabant 6:3 et cla-

is the LORD of hosts: the whole earth is full of his glory.

mabant alter ad alterum et dicebant sanctus sanctus sanctus

6:4 And the posts of the door moved at the voice of him Dominus exercituum plena est omnis terra gloria eius 6:4

that cried, and the house was filled with smoke.

et commota sunt superliminaria cardinum a uoce clamantis

6:5 Then said I, Woe is me! for I am undone; because I et domus impleta est fumo 6:5 et dixi uae mihi quia tacui

am a man of unclean lips, and I dwell in the midst of a people of unclean lips: for mine eyes have seen the King,

the LORD of hosts. quia uir pollutus labiis ego sum et in medio populi polluta

labia habentis ego habito et Regem Dominum exercituum

6:6 Then flew one of the seraphims unto me, having a live uidi oculis meis 6:6 et uolauit ad me unus de seraphin et

coal in his hand, which he had taken with the tongs from off the altar:

in manu eius calculus quem forcipe tulerat de altari 6:7 et

6:7 And he laid it upon my mouth, and said, Lo, this hath touched thy lips; and thine iniquity is taken away, and thy sin purged.

tetigit os meum et dixit ecce tetigit hoc labia tua et aufere-

6:8 Also I heard the voice of the Lord, saying, Whom shall tur iniquitas tua et peccatum tuum mundabitur 6:8 et au-I send, and who will go for us? Then said I, Here am I; send me.

diui uocem Domini dicentis quem mittam et quis ibit nobis

6:9 And he said, Go, and tell this people, Hear ye indeed, et dixi ecce ego sum mitte me 6:9 et dixit uade et dices pobut understand not; and see ye indeed, but perceive not.

pulo huic audite audientes et nolite intellegere et uidete ui-

6:10 Make the heart of this people fat, and make their ears sionem et nolite cognoscere 6:10 excaeca cor populi huius

heavy, and shut their eyes; lest they see with their eyes, and hear with their ears, and understand with their heart,

and convert, and be healed. et aures eius adgraua et oculos eius claude ne forte uideat

oculis suis et auribus suis audiat et corde suo intellegat et

6:11 Then said I, Lord, how long? And he answered, Un- conuertatur et sanem eum 6:11 et dixi usquequo Domine et

til the cities be wasted without inhabitant, and the houses without man, and the land be utterly desolate,

dixit donec desolentur ciuitates absque habitatore et domus

6:12 And the LORD have removed men far away, and sine homine et terra relinquetur deserta 6:12 et longe faciet

there be a great forsaking in the midst of the land.

Dominus homines et multiplicabitur quae derelicta fuerat in

6:13 But yet in it shall be a tenth, and it shall return, and medio terrae 6:13 et adhuc in ea decimatio et conuertetur et

shall be eaten: as a teil tree, and as an oak, whose substance is in them, when they cast their leaves: so the holy

seed shall be the substance thereof. erit in ostensionem sicut terebinthus et sicuti quercus quae expandit ramos suos semen sanctum erit id quod steterit in ea

7:1 And it came to pass in the days of Ahaz the son of 7:1 et factum est in diebus Ahaz filii Ioatham filii Oziae

Jotham, the son of Uzziah, king of Judah, that Rezin the king of Syria, and Pekah the son of Remaliah, king of Israel, went up toward Jerusalem to war against it, but could not prevail against it.

regis Iuda ascendit Rasin rex Syriae et Phacee filius Romeliae rex Israhel in Hierusalem ad proeliandum contra eam

7:2 And it was told the house of David, saying, Syria is et non potuerunt debellare eam 7:2 et nuntiauerunt domui

confederate with Ephraim. And his heart was moved, and the heart of his people, as the trees of the wood are moved

with the wind. Dauid dicentes requieuit Syria super Ephraim et commotum

est cor eius et cor populi eius sicut mouentur ligna siluarum

7:3 Then said the LORD unto Isaiah, Go forth now to meet a facie uenti 7:3 et dixit Dominus ad Isaiam egredere in

Ahaz, thou, and Shearjashub thy son, at the end of the conduit of the upper pool in the highway of the fuller’s field;

Prophetia Isaiae

occursum Ahaz tu et qui derelictus est Iasub filius tuus ad extremum aquaeductus piscinae superioris in uia agri Fullonis 7:4 et dices ad eum uide ut sileas noli timere et cor 7:4 And say unto him, Take heed, and be quiet; fear not,

neither be fainthearted for the two tails of these smoking firebrands, for the fierce anger of Rezin with Syria, and of

tuum ne formidet a duobus caudis titionum fumigantium is- the son of Remaliah. torum in ira furoris Rasin et Syriae et filii Romeliae 7:5 7:5 Because Syria, Ephraim, and the son of Remaliah,

have taken evil counsel against thee, saying,

eo quod consilium inierit contra te Syria malum Ephraim et filius Romeliae dicentes 7:6 ascendamus ad Iudam et sus- 7:6 Let us go up against Judah, and vex it, and let us make

a breach therein for us, and set a king in the midst of it, even the son of Tabeal:

citemus eum et auellamus eum ad nos et ponamus regem in medio eius filium Tabeel 7:7 haec dicit Dominus Deus non 7:7 Thus saith the Lord GOD, It shall not stand, neither

shall it come to pass.

stabit et non erit istud 7:8 sed caput Syriae Damascus et 7:8 For the head of Syria is Damascus, and the head of

Damascus is Rezin; and within threescore and five years shall Ephraim be broken, that it be not a people.

caput Damasci Rasin et adhuc sexaginta et quinque anni et desinet Ephraim esse populus 7:9 et caput Ephraim Sama- 7:9 And the head of Ephraim is Samaria, and the head of

Samaria is Remaliah’s son. If ye will not believe, surely ye shall not be established.

ria et caput Samariae filius Romeliae si non credideritis non permanebitis 7:10 et adiecit Dominus loqui ad Ahaz dicens 7:10 Moreover the LORD spake again unto Ahaz, saying, 7:11 pete tibi signum a Domino Deo tuo in profundum in- 7:11 Ask thee a sign of the LORD thy God; ask it either in

the depth, or in the height above.

ferni siue in excelsum supra 7:12 et dixit Ahaz non petam 7:12 But Ahaz said, I will not ask, neither will I tempt the

LORD.

et non temptabo Dominum 7:13 et dixit audite ergo domus 7:13 And he said, Hear ye now, O house of David; Is it a

small thing for you to weary men, but will ye weary my God also?

Dauid numquid parum uobis est molestos esse hominibus quia molesti estis et Deo meo 7:14 propter hoc dabit Domi- 7:14 Therefore the Lord himself shall give you a sign; Be-

hold, a virgin shall conceive, and bear a son, and shall call his name Immanuel.

nus ipse uobis signum ecce uirgo concipiet et pariet filium et uocabitis nomen eius Emmanuhel 7:15 butyrum et mel co- 7:15 Butter and honey shall he eat, that he may know to

refuse the evil, and choose the good.

medet ut sciat reprobare malum et eligere bonum 7:16 quia 7:16 For before the child shall know to refuse the evil,

and choose the good, the land that thou abhorrest shall be forsaken of both her kings.

antequam sciat puer reprobare malum et eligere bonum derelinquetur terra quam tu detestaris a facie duum regum suorum 7:17 adducet Dominus super te et super populum tuum 7:17 The LORD shall bring upon thee, and upon thy peo-

ple, and upon thy father’s house, days that have not come, from the day that Ephraim departed from Judah; even the

et super domum patris tui dies qui non uenerunt a diebus king of Assyria. separationis Ephraim a Iuda cum rege Assyriorum 7:18 et 7:18 And it shall come to pass in that day, that the LORD

shall hiss for the fly that is in the uttermost part of the rivers of Egypt, and for the bee that is in the land of As-

erit in die illa sibilabit Dominus muscae quae est in extremo syria. fluminum Aegypti et api quae est in terra Assur 7:19 et ue- 7:19 And they shall come, and shall rest all of them in the

desolate valleys, and in the holes of the rocks, and upon all thorns, and upon all bushes.

nient et requiescent omnes in torrentibus uallium et cauernis petrarum et in omnibus frutectis et in uniuersis foraminibus 7:20 in die illa radet Dominus in nouacula conducta in his qui 7:20 In the same day shall the Lord shave with a razor that

is hired, namely, by them beyond the river, by the king of Assyria, the head, and the hair of the feet: and it shall also

trans Flumen sunt in rege Assyriorum caput et pilos pedum consume the beard.

Biblia Sacra Vulgata

7:21 And it shall come to pass in that day, that a man shall et barbam uniuersam 7:21 et erit in die illa nutriet homo nourish a young cow, and two sheep;

7:22 And it shall come to pass, for the abundance of milk uaccam boum et duas oues 7:22 et prae ubertate lactis co-

that they shall give he shall eat butter: for butter and honey shall every one eat that is left in the land.

medet butyrum butyrum enim et mel manducabit omnis qui

7:23 And it shall come to pass in that day, that every place relictus fuerit in medio terrae 7:23 et erit in die illa omnis

shall be, where there were a thousand vines at a thousand silverlings, it shall even be for briers and thorns.

locus ubi fuerint mille uites mille argenteis et in spinas et

7:24 With arrows and with bows shall men come thither; in uepres erunt 7:24 cum sagittis et arcu ingredientur illuc

because all the land shall become briers and thorns. 7:25 And on all hills that shall be digged with the mattock, uepres enim et spinae erunt in uniuersa terra 7:25 et omnes there shall not come thither the fear of briers and thorns: but it shall be for the sending forth of oxen, and for the

treading of lesser cattle. montes qui in sarculo sarientur non ueniet illuc terror spinarum et ueprium et erit in pascua bouis et in conculcationem pecoris

8:1 Moreover the LORD said unto me, Take thee a great 8:1 et dixit Dominus ad me sume tibi librum grandem et

roll, and write in it with a man’s pen concerning Mahershalalhashbaz.

scribe in eo stilo hominis Velociter spolia detrahe Cito prae-

8:2 And I took unto me faithful witnesses to record, Uriah dare 8:2 et adhibui mihi testes fideles Vriam sacerdotem the priest, and Zechariah the son of Jeberechiah.

8:3 And I went unto the prophetess; and she conceived, et Zacchariam filium Barachiae 8:3 et accessi ad prophe-

and bare a son. Then said the LORD to me, Call his name Mahershalalhashbaz.

tissam et concepit et peperit filium et dixit Dominus ad me uoca nomen eius Adcelera spolia detrahere Festina praedari

8:4 For before the child shall have knowledge to cry, My 8:4 quia antequam sciat puer uocare patrem suum et matrem

father, and my mother, the riches of Damascus and the spoil of Samaria shall be taken away before the king of

Assyria. suam auferetur fortitudo Damasci et spolia Samariae coram 8:5 The LORD spake also unto me again, saying, rege Assyriorum 8:5 et adiecit Dominus loqui ad me adhuc 8:6 Forasmuch as this people refuseth the waters of dicens 8:6 pro eo quod abiecit populus iste aquas Siloae

Shiloah that go softly, and rejoice in Rezin and Remaliah’s son;

quae uadunt cum silentio et adsumpsit magis Rasin et filium

8:7 Now therefore, behold, the Lord bringeth up upon Romeliae 8:7 propter hoc ecce Dominus adducet super eos

them the waters of the river, strong and many, even the king of Assyria, and all his glory: and he shall come up

over all his channels, and go over all his banks: aquas Fluminis fortes et multas regem Assyriorum et om-

nem gloriam eius et ascendet super omnes riuos eius et fluet

8:8 And he shall pass through Judah; he shall overflow and super uniuersas ripas eius 8:8 et ibit per Iudam inundans et

go over, he shall reach even to the neck; and the stretching out of his wings shall fill the breadth of thy land, O

Immanuel. transiens usque ad collum ueniet et erit extensio alarum eius 8:9 Associate yourselves, O ye people, and ye shall be implens latitudinem terrae tuae o Emmanuhel 8:9 congre-

broken in pieces; and give ear, all ye of far countries: gird yourselves, and ye shall be broken in pieces; gird your-

selves, and ye shall be broken in pieces. gamini populi et uincimini et audite uniuersae procul terrae 8:10 Take counsel together, and it shall come to nought; confortamini et uincimini accingite uos et uincimini 8:10

speak the word, and it shall not stand: for God is with us.

inite consilium et dissipabitur loquimini uerbum et non fiet

8:11 For the LORD spake thus to me with a strong hand, quia nobiscum Deus 8:11 haec enim ait Dominus ad me and instructed me that I should not walk in the way of this people, saying,

sicut in forti manu erudiuit me ne irem in uia populi huius

8:12 Say ye not, A confederacy, to all them to whom this dicens 8:12 non dicatis coniuratio omnia enim quae loqui-

people shall say, A confederacy; neither fear ye their fear, nor be afraid.

Prophetia Isaiae

tur populus iste coniuratio est et timorem eius ne timeatis neque paueatis 8:13 Dominum exercituum ipsum sanctifi- 8:13 Sanctify the LORD of hosts himself; and let him be

your fear, and let him be your dread.

cate ipse pauor uester et ipse terror uester 8:14 et erit uobis 8:14 And he shall be for a sanctuary; but for a stone of

stumbling and for a rock of offence to both the houses of Israel, for a gin and for a snare to the inhabitants of

in sanctificationem in lapidem autem offensionis et in pet- Jerusalem. ram scandali duabus domibus Israhel in laqueum et in ruinam habitantibus Hierusalem 8:15 et offendent ex eis plu- 8:15 And many among them shall stumble, and fall, and

be broken, and be snared, and be taken.

rimi et cadent et conterentur et inretientur et capientur 8:16 8:16 Bind up the testimony, seal the law among my disci-

ples.

liga testimonium signa legem in discipulis meis 8:17 et ex- 8:17 And I will wait upon the LORD, that hideth his face

from the house of Jacob, and I will look for him.

pectabo Dominum qui abscondit faciem suam a domo Iacob et praestolabor eum 8:18 ecce ego et pueri quos mihi de- 8:18 Behold, I and the children whom the LORD hath

given me are for signs and for wonders in Israel from the LORD of hosts, which dwelleth in mount Zion.

dit Dominus in signum et in portentum Israhelis a Domino exercituum qui habitat in monte Sion 8:19 et cum dixerint 8:19 And when they shall say unto you, Seek unto them

that have familiar spirits, and unto wizards that peep, and that mutter: should not a people seek unto their God? for

ad uos quaerite a pythonibus et a diuinis qui stridunt in in- the living to the dead? cantationibus suis numquid non populus a Deo suo requirit pro uiuis a mortuis 8:20 ad legem magis et ad testimonium 8:20 To the law and to the testimony: if they speak not

according to this word, it is because there is no light in them.

quod si non dixerint iuxta uerbum hoc non erit eis matutina lux 8:21 et transibit per eam corruet et esuriet et cum esu- 8:21 And they shall pass through it, hardly bestead and

hungry: and it shall come to pass, that when they shall be hungry, they shall fret themselves, and curse their king and

rierit irascetur et maledicet regi suo et Deo suo et suspiciet their God, and look upward. sursum 8:22 et ad terram intuebitur et ecce tribulatio et te- 8:22 And they shall look unto the earth; and behold trouble

and darkness, dimness of anguish; and they shall be driven to darkness.

nebrae dissolutio angustia et caligo persequens et non poterit auolare de angustia sua

9:1 primo tempore adleuiata est terra Zabulon et terra Nep- 9:1 Nevertheless the dimness shall not be such as was in

her vexation, when at the first he lightly afflicted the land of Zebulun and the land of Naphtali, and afterward did more grievously afflict her by the way of the sea, beyond Jordan, in Galilee of the nations.

thalim et nouissimo adgrauata est uia maris trans Iordanem Galileae gentium 9:2 populus qui ambulabat in tenebris ui- 9:2 The people that walked in darkness have seen a great

light: they that dwell in the land of the shadow of death, upon them hath the light shined.

dit lucem magnam habitantibus in regione umbrae mortis lux orta est eis 9:3 multiplicasti gentem non magnificasti laeti- 9:3 Thou hast multiplied the nation, and not increased the

joy: they joy before thee according to the joy in harvest, and as men rejoice when they divide the spoil.

tiam laetabuntur coram te sicut laetantur in messe sicut exultant quando diuidunt spolia 9:4 iugum enim oneris eius et 9:4 For thou hast broken the yoke of his burden, and the

staff of his shoulder, the rod of his oppressor, as in the day of Midian.

uirgam umeri eius et sceptrum exactoris eius superasti sicut

in die Madian 9:5 quia omnis uiolenta praedatio cum tu- 9:5 For every battle of the warrior is with confused noise, and garments rolled in blood; but this shall be with burning and fuel of fire.

multu et uestimentum mixtum sanguine erit in conbustionem et cibus ignis 9:6 paruulus enim natus est nobis filius 9:6 For unto us a child is born, unto us a son is given: and

the government shall be upon his shoulder: and his name shall be called Wonderful, Counsellor, The mighty God,

datus est nobis et factus est principatus super umerum eius The everlasting Father, The Prince of Peace.

Biblia Sacra Vulgata

et uocabitur nomen eius Admirabilis consiliarius Deus fortis

9:7 Of the increase of his government and peace there shall Pater futuri saeculi Princeps pacis 9:7 multiplicabitur eius

be no end, upon the throne of David, and upon his kingdom, to order it, and to establish it with judgment and with justice from henceforth even for ever. The zeal of the LORD of hosts will perform this.

imperium et pacis non erit finis super solium Dauid et super regnum eius ut confirmet illud et corroboret in iudicio et iustitia amodo et usque in sempiternum zelus Domini ex-

9:8 The Lord sent a word into Jacob, and it hath lighted ercituum faciet hoc 9:8 uerbum misit Dominus in Iacob et

upon Israel.

9:9 And all the people shall know, even Ephraim and the cecidit in Israhel 9:9 et sciet populus omnis Ephraim et ha-

inhabitant of Samaria, that say in the pride and stoutness of heart,

bitantes Samariam in superbia et magnitudine cordis dicen-

9:10 The bricks are fallen down, but we will build with tes 9:10 lateres ceciderunt sed quadris lapidibus aedificabi-

hewn stones: the sycomores are cut down, but we will change them into cedars.

mus sycomoros succiderunt sed cedros inmutabimus 9:11

9:11 Therefore the LORD shall set up the adversaries of Rezin against him, and join his enemies together;

et eleuabit Dominus hostes Rasin super eum et inimicos eius

9:12 The Syrians before, and the Philistines behind; and in tumultum uertet 9:12 Syriam ab oriente et Philisthim ab

they shall devour Israel with open mouth. For all this his anger is not turned away, but his hand is stretched out still.

occidente et deuorabunt Israhel toto ore in omnibus his non

9:13 For the people turneth not unto him that smiteth them, est auersus furor eius sed adhuc manus eius extenta 9:13 neither do they seek the LORD of hosts.

et populus non est reuersus ad percutientem se et Dominum

9:14 Therefore the LORD will cut off from Israel head and exercituum non inquisierunt 9:14 et disperdet Dominus ab

tail, branch and rush, in one day.

Israhel caput et caudam incuruantem et refrenantem die una

9:15 The ancient and honourable, he is the head; and the 9:15 longeuus et honorabilis ipse est caput et propheta do-

prophet that teacheth lies, he is the tail.

9:16 For the leaders of this people cause them to err; and cens mendacium ipse cauda est 9:16 et erunt qui beatificant

they that are led of them are destroyed.

populum istum seducentes et qui beatificantur praecipitati

9:17 Therefore the LORD shall have no joy in their young 9:17 propter hoc super adulescentulis eius non laetabitur Do-

men, neither shall have mercy on their fatherless and widows: for every one is an hypocrite and an evildoer, and every mouth speaketh folly. For all this his anger is not turned away, but his hand is stretched out still.

minus et pupillorum eius et uiduarum non miserebitur quia omnis hypocrita est et nequam et uniuersum os locutum est stultitiam in omnibus his non est auersus furor eius sed ad-

9:18 For wickedness burneth as the fire: it shall devour huc manus eius extenta 9:18 succensa est enim quasi igthe briers and thorns, and shall kindle in the thickets of the forest, and they shall mount up like the lifting up of

smoke. nis impietas ueprem et spinam uorabit et succendetur in de-9:19 Through the wrath of the LORD of hosts is the land nsitate saltus et conuoluetur superbia fumi 9:19 in ira Do-

darkened, and the people shall be as the fuel of the fire: no man shall spare his brother.

mini exercituum conturbata est terra et erit populus quasi

9:20 And he shall snatch on the right hand, and be hungry; esca ignis uir fratri suo non parcet 9:20 et declinabit ad de-

and he shall eat on the left hand, and they shall not be satisfied: they shall eat every man the flesh of his own

arm: xteram et esuriet et comedet ad sinistram et non saturabitur

unusquisque carnem brachii sui uorabit Manasses Ephraim

9:21 Manasseh, Ephraim; and Ephraim, Manasseh: and et Ephraim Manassen simul ipsi contra Iudam 9:21 in om-

they together shall be against Judah. For all this his anger is not turned away, but his hand is stretched out still.

nibus his non est auersus furor eius sed adhuc manus eius

Prophetia Isaiae

extenta

10:1 uae qui condunt leges iniquas et scribentes iniustitiam 10:1 Woe unto them that decree unrighteous decrees, and

that write grievousness which they have prescribed;

scripserunt 10:2 ut opprimerent in iudicio pauperes et uim 10:2 To turn aside the needy from judgment, and to take

away the right from the poor of my people, that widows may be their prey, and that they may rob the fatherless!

facerent causae humilium populi mei ut essent uiduae praeda eorum et pupillos diriperent 10:3 quid facietis in die ui- 10:3 And what will ye do in the day of visitation, and in

the desolation which shall come from far? to whom will ye flee for help? and where will ye leave your glory?

sitationis et calamitatis de longe uenientis ad cuius fugietis auxilium et ubi derelinquetis gloriam uestram 10:4 ne in- 10:4 Without me they shall bow down under the prisoners,

and they shall fall under the slain. For all this his anger is not turned away, but his hand is stretched out still.

curuemini sub uinculo et cum interfectis cadatis super omnibus his non est auersus furor eius sed adhuc manus eius extenta 10:5 uae Assur uirga furoris mei et baculus ipse in 10:5 O Assyrian, the rod of mine anger, and the staff in

their hand is mine indignation.

manu eorum indignatio mea 10:6 ad gentem fallacem mit- 10:6 I will send him against an hypocritical nation, and against the people of my wrath will I give him a charge, to take the spoil, and to take the prey, and to tread them

tam eum et contra populum furoris mei mandabo illi ut aufe- down like the mire of the streets. rat spolia et diripiat praedam et ponat illum in conculcationem quasi lutum platearum 10:7 ipse autem non sic arbit- 10:7 Howbeit he meaneth not so, neither doth his heart

think so; but it is in his heart to destroy and cut off nations not a few.

rabitur et cor eius non ita aestimabit sed ad conterendum erit cor eius et ad internicionem gentium non paucarum 10:8 di- 10:8 For he saith, Are not my princes altogether kings? cet enim 10:9 numquid non principes mei simul reges sunt 10:9 Is not Calno as Carchemish? is not Hamath as Arpad?

is not Samaria as Damascus?

numquid non ut Charchamis sic Chalanno et ut Arfad sic Emath numquid non ut Damascus sic Samaria 10:10 quo- 10:10 As my hand hath found the kingdoms of the idols,

and whose graven images did excel them of Jerusalem and of Samaria;

modo inuenit manus mea regna idoli sic et simulacra eorum de Hierusalem et de Samaria 10:11 numquid non sic- 10:11 Shall I not, as I have done unto Samaria and her

idols, so do to Jerusalem and her idols?

ut feci Samariae et idolis eius sic faciam Hierusalem et simulacris eius 10:12 et erit cum impleuerit Dominus cuncta 10:12 Wherefore it shall come to pass, that when the Lord

hath performed his whole work upon mount Zion and on Jerusalem, I will punish the fruit of the stout heart of the

opera sua in monte Sion et in Hierusalem uisitabo super fruc- king of Assyria, and the glory of his high looks. tum magnifici cordis regis Assur et super gloriam altitudinis oculorum eius 10:13 dixit enim in fortitudine manus meae 10:13 For he saith, By the strength of my hand I have done

it, and by my wisdom; for I am prudent: and I have removed the bounds of the people, and have robbed their treasures, and I have put down the inhabitants like a valiant man:

feci et in sapientia mea intellexi et abstuli terminos populorum et principes eorum depraedatus sum et detraxi quasi potens in sublime residentes 10:14 et inuenit quasi nidum 10:14 And my hand hath found as a nest the riches of the

people: and as one gathereth eggs that are left, have I gathered all the earth; and there was none that moved the wing,

manus mea fortitudinem populorum et sicut colliguntur oua or opened the mouth, or peeped. quae derelicta sunt sic uniuersam terram ego congregaui et non fuit qui moueret pinnam et aperiret os et ganniret 10:15 10:15 Shall the axe boast itself against him that heweth

therewith? or shall the saw magnify itself against him that shaketh it? as if the rod should shake itself against them that lift it up, or as if the staff should lift up itself, as if it were no wood.

numquid gloriabitur securis contra eum qui secat in ea aut

Biblia Sacra Vulgata

exaltabitur serra contra eum a quo trahitur quomodo si eleuetur uirga contra leuantem se et exaltetur baculus qui uti-

10:16 Therefore shall the Lord, the Lord of hosts, send que lignum est 10:16 propter hoc mittet Dominator Deus among his fat ones leanness; and under his glory he shall kindle a burning like the burning of a fire.

exercituum in pinguibus eius tenuitatem et subtus gloriam

10:17 And the light of Israel shall be for a fire, and his eius succensa ardebit quasi conbustio ignis 10:17 et erit lu-

Holy One for a flame: and it shall burn and devour his thorns and his briers in one day;

men Israhel in igne et Sanctus eius in flamma et succende-

10:18 And shall consume the glory of his forest, and of tur et deuorabitur spina eius et uepres in die una 10:18 et

his fruitful field, both soul and body: and they shall be as when a standard-bearer fainteth.

gloria saltus eius et Carmeli eius ab anima usque ad car-

10:19 And the rest of the trees of his forest shall be few, nem consumetur et erit terrore profugus 10:19 et reliquiae

that a child may write them.

ligni saltus eius pro paucitate numerabuntur et puer scribet

10:20 And it shall come to pass in that day, that the rem- eos 10:20 et erit in die illa non adiciet residuum Israhel et

nant of Israel, and such as are escaped of the house of Jacob, shall no more again stay upon him that smote them; but shall stay upon the LORD, the Holy One of Israel, in truth.

hii qui fugerint de domo Iacob inniti super eo qui percutit eos sed innitetur super Dominum Sanctum Israhel in ue-

10:21 The remnant shall return, even the remnant of Jacob, ritate 10:21 reliquiae conuertentur reliquiae inquam Iacob

unto the mighty God.

10:22 For though thy people Israel be as the sand of the ad Deum fortem 10:22 si enim fuerit populus tuus Israhel

sea, yet a remnant of them shall return: the consumption decreed shall overflow with righteousness.

quasi harena maris reliquiae conuertentur ex eo consumma-

10:23 For the Lord GOD of hosts shall make a consump- tio adbreuiata inundabit iustitiam 10:23 consummationem

tion, even determined, in the midst of all the land.

enim et adbreuiationem Dominus Deus exercituum faciet in

10:24 Therefore thus saith the Lord GOD of hosts, O my medio omnis terrae 10:24 propter hoc haec dicit Dominus

people that dwellest in Zion, be not afraid of the Assyrian: he shall smite thee with a rod, and shall lift up his staff

against thee, after the manner of Egypt. Deus exercituum noli timere populus meus habitator Sion

ab Assur in uirga percutiet te et baculum suum leuabit super

10:25 For yet a very little while, and the indignation shall te in uia Aegypti 10:25 adhuc enim paululum modicumque cease, and mine anger in their destruction.

et consummabitur indignatio et furor meus super scelus eo-

10:26 And the LORD of hosts shall stir up a scourge for rum 10:26 et suscitabit super eum Dominus exercituum fla-

him according to the slaughter of Midian at the rock of Oreb: and as his rod was upon the sea, so shall he lift it up

after the manner of Egypt. gellum iuxta plagam Madian in petra Oreb et uirgam suam

10:27 And it shall come to pass in that day, that his burden super mare et leuabit eam in uia Aegypti 10:27 et erit in shall be taken away from off thy shoulder, and his yoke from off thy neck, and the yoke shall be destroyed because

of the anointing. die illa auferetur onus eius de umero tuo et iugum eius de 10:28 He is come to Aiath, he is passed to Migron; at collo tuo et conputrescet iugum a facie olei 10:28 ueniet in

Michmash he hath laid up his carriages:

Aiath transibit in Magron apud Machmas commendabit uasa

10:29 They are gone over the passage: they have taken up sua 10:29 transierunt cursim Gabee sedes nostra obstipuit

their lodging at Geba; Ramah is afraid; Gibeah of Saul is fled.

Rama Gabaath Saulis fugit 10:30 hinni uoce tua filia Gal-

10:30 Lift up thy voice, O daughter of Gallim: cause it to be heard unto Laish, O poor Anathoth.

10:31 Madmenah is removed; the inhabitants of Gebim lim adtende Laisa paupercula Anathoth 10:31 migrauit Me-

gather themselves to flee.

10:32 As yet shall he remain at Nob that day: he shall demena habitatores Gebim confortamini 10:32 adhuc dies

shake his hand against the mount of the daughter of Zion, the hill of Jerusalem.

Prophetia Isaiae

est ut in Nob stetur agitabit manum suam super montem filiae Sion collem Hierusalem 10:33 ecce Dominator Domi- 10:33 Behold, the Lord, the LORD of hosts, shall lop the

bough with terror: and the high ones of stature shall be hewn down, and the haughty shall be humbled.

nus exercituum confringet lagunculam in terrore et excelsi statura succidentur et sublimes humiliabuntur 10:34 et sub- 10:34 And he shall cut down the thickets of the forest with

iron, and Lebanon shall fall by a mighty one.

uertentur condensa saltus ferro et Libanus cum excelsis cadet

11:1 et egredietur uirga de radice Iesse et flos de radice 11:1 And there shall come forth a rod out of the stem of

Jesse, and a Branch shall grow out of his roots:

eius ascendet 11:2 et requiescet super eum spiritus Domini 11:2 And the spirit of the LORD shall rest upon him, the

spirit of wisdom and understanding, the spirit of counsel and might, the spirit of knowledge and of the fear of the

spiritus sapientiae et intellectus spiritus consilii et fortitudi- LORD; nis spiritus scientiae et pietatis 11:3 et replebit eum spiritus 11:3 And shall make him of quick understanding in the

fear of the LORD: and he shall not judge after the sight of his eyes, neither reprove after the hearing of his ears:

timoris Domini non secundum uisionem oculorum iudicabit neque secundum auditum aurium arguet 11:4 sed iudicabit 11:4 But with righteousness shall he judge the poor, and

reprove with equity for the meek of the earth: and he shall smite the earth: with the rod of his mouth, and with the

in iustitia pauperes et arguet in aequitate pro mansuetis ter- breath of his lips shall he slay the wicked. rae et percutiet terram uirga oris sui et spiritu labiorum suo-

rum interficiet impium 11:5 et erit iustitia cingulum lum- 11:5 And righteousness shall be the girdle of his loins, and

faithfulness the girdle of his reins.

borum eius et fides cinctorium renis eius 11:6 habitabit lu- 11:6 The wolf also shall dwell with the lamb, and the leop-

ard shall lie down with the kid; and the calf and the young lion and the fatling together; and a little child shall lead

pus cum agno et pardus cum hedo accubabit uitulus et leo them. et ouis simul morabuntur et puer paruulus minabit eos 11:7 11:7 And the cow and the bear shall feed; their young ones

shall lie down together: and the lion shall eat straw like the ox.

uitulus et ursus pascentur simul requiescent catuli eorum et leo quasi bos comedet paleas 11:8 et delectabitur infans ab 11:8 And the sucking child shall play on the hole of the

asp, and the weaned child shall put his hand on the cockatrice’ den.

ubere super foramine aspidis et in cauerna reguli qui ablactatus fuerit manum suam mittet 11:9 non nocebunt et non 11:9 They shall not hurt nor destroy in all my holy moun-

tain: for the earth shall be full of the knowledge of the LORD, as the waters cover the sea.

occident in uniuerso monte sancto meo quia repleta est terra scientia Domini sicut aquae maris operientes 11:10 in die 11:10 And in that day there shall be a root of Jesse, which

shall stand for an ensign of the people; to it shall the Gentiles seek: and his rest shall be glorious.

illa radix Iesse qui stat in signum populorum ipsum gentes deprecabuntur et erit sepulchrum eius gloriosum 11:11 et 11:11 And it shall come to pass in that day, that the Lord

shall set his hand again the second time to recover the remnant of his people, which shall be left, from Assyria, and from Egypt, and from Pathros, and from Cush, and from Elam, and from Shinar, and from Hamath, and from the islands of the sea.

erit in die illa adiciet Dominus secundo manum suam ad possidendum residuum populi sui quod relinquetur ab Assyriis et ab Aegypto et a Fetros et ab Aethiopia et ab Aelam et a Sennaar et ab Emath et ab insulis maris 11:12 et leuabit 11:12 And he shall set up an ensign for the nations, and

shall assemble the outcasts of Israel, and gather together the dispersed of Judah from the four corners of the earth.

signum in nationes et congregabit profugos Israhel et dispersos Iuda colliget a quattuor plagis terrae 11:13 et auferetur 11:13 The envy also of Ephraim shall depart, and the ad-

versaries of Judah shall be cut off: Ephraim shall not envy Judah, and Judah shall not vex Ephraim.

zelus Ephraim et hostes Iuda peribunt Ephraim non aemu-

Biblia Sacra Vulgata

11:14 But they shall fly upon the shoulders of the labitur Iudam et Iudas non pugnabit contra Ephraim 11:14

Philistines toward the west; they shall spoil them of the east together: they shall lay their hand upon Edom and

Moab; and the children of Ammon shall obey them. et uolabunt in umeros Philisthim per mare simul praedabun-

tur filios orientis Idumea et Moab praeceptum manus eorum

11:15 And the LORD shall utterly destroy the tongue of et filii Ammon oboedientes erunt 11:15 et desolabit Domi-

the Egyptian sea; and with his mighty wind shall he shake his hand over the river, and shall smite it in the seven

streams, and make men go over dryshod. nus linguam maris Aegypti et leuabit manum suam super

Flumen in fortitudine spiritus sui et percutiet eum in septem

11:16 And there shall be an highway for the remnant of his riuis ita ut transeant per eum calciati 11:16 et erit uia resi-

people, which shall be left, from Assyria; like as it was to Israel in the day that he came up out of the land of Egypt.

duo populo meo qui relinquetur ab Assyriis sicut fuit Israhel in die qua ascendit de terra Aegypti

12:1 And in that day thou shalt say, O LORD, I will praise 12:1 et dices in illa die confitebor tibi Domine quoniam thee: though thou wast angry with me, thine anger is turned away, and thou comfortedst me.

iratus es mihi conuersus est furor tuus et consolatus es me

12:2 Behold, God is my salvation; I will trust, and not be 12:2 ecce Deus saluator meus fiducialiter agam et non tiafraid: for the LORD JEHOVAH is my strength and my song; he also is become my salvation.

mebo quia fortitudo mea et laus mea Dominus Deus et factus

12:3 Therefore with joy shall ye draw water out of the est mihi in salutem 12:3 haurietis aquas in gaudio de fonti-

wells of salvation.

12:4 And in that day shall ye say, Praise the LORD, call bus saluatoris 12:4 et dicetis in illa die confitemini Domino

upon his name, declare his doings among the people, make mention that his name is exalted.

et inuocate nomen eius notas facite in populis adinuentio-

12:5 Sing unto the LORD; for he hath done excellent nes eius mementote quoniam excelsum est nomen eius 12:5

things: this is known in all the earth.

cantate Domino quoniam magnifice fecit adnuntiate hoc in

12:6 Cry out and shout, thou inhabitant of Zion: for great uniuersa terra 12:6 exulta et lauda habitatio Sion quia ma-

is the Holy One of Israel in the midst of thee.

gnus in medio tui Sanctus Israhel

13:1 The burden of Babylon, which Isaiah the son of 13:1 onus Babylonis quod uidit Isaias filius Amos 13:2

Amoz did see. 13:2 Lift ye up a banner upon the high mountain, exalt the voice unto them, shake the hand, that they may go into the gates of the nobles.

super montem caligosum leuate signum exaltate uocem le-

13:3 I have commanded my sanctified ones, I have also uate manum et ingrediantur portas duces 13:3 ego mandaui

called my mighty ones for mine anger, even them that rejoice in my highness.

sanctificatis meis et uocaui fortes meos in ira mea exultantes

13:4 The noise of a multitude in the mountains, like as in gloria mea 13:4 uox multitudinis in montibus quasi po-

of a great people; a tumultuous noise of the kingdoms of nations gathered together: the LORD of hosts mustereth

the host of the battle. pulorum frequentium uox sonitus regum gentium congrega-13:5 They come from a far country, from the end of tarum Dominus exercituum praecepit militiae belli 13:5 ue-

heaven, even the LORD, and the weapons of his indignation, to destroy the whole land.

nientibus de terra procul a summitate caeli Dominus et uasa

13:6 Howl ye; for the day of the LORD is at hand; it shall furoris eius ut disperdat omnem terram 13:6 ululate quia

come as a destruction from the Almighty.

13:7 Therefore shall all hands be faint, and every man’s prope est dies Domini quasi uastitas a Domino ueniet 13:7

heart shall melt:

propter hoc omnes manus dissoluentur et omne cor homi-

13:8 And they shall be afraid: pangs and sorrows shall nis tabescet 13:8 et conteretur tortiones et dolores tenebunt

take hold of them; they shall be in pain as a woman that travaileth: they shall be amazed one at another; their faces

shall be as flames. quasi parturiens dolebunt unusquisque ad proximum suum

Prophetia Isaiae

stupebit facies conbustae uultus eorum 13:9 ecce dies Do- 13:9 Behold, the day of the LORD cometh, cruel both with

wrath and fierce anger, to lay the land desolate: and he shall destroy the sinners thereof out of it.

mini uenit crudelis et indignationis plenus et irae furorisque ad ponendam terram in solitudine et peccatores eius conte-

rendos de ea 13:10 quoniam stellae caeli et splendor earum 13:10 For the stars of heaven and the constellations thereof shall not give their light: the sun shall be darkened in his going forth, and the moon shall not cause her light to

non expandent lumen suum obtenebratus est sol in ortu suo shine. et luna non splendebit in lumine suo 13:11 et uisitabo su- 13:11 And I will punish the world for their evil, and the

wicked for their iniquity; and I will cause the arrogancy of the proud to cease, and will lay low the haughtiness of the

per orbis mala et contra impios iniquitatem eorum et quie- terrible. scere faciam superbiam infidelium et arrogantiam fortium humiliabo 13:12 pretiosior erit uir auro et homo mundo ob- 13:12 I will make a man more precious than fine gold;

even a man than the golden wedge of Ophir.

rizo 13:13 super hoc caelum turbabo et mouebitur terra de 13:13 Therefore I will shake the heavens, and the earth

shall remove out of her place, in the wrath of the LORD of hosts, and in the day of his fierce anger.

loco suo propter indignationem Domini exercituum et propter diem irae furoris eius 13:14 et erit quasi dammula fu- 13:14 And it shall be as the chased roe, and as a sheep that

no man taketh up: they shall every man turn to his own people, and flee every one into his own land.

giens et quasi ouis et non erit qui congreget unusquisque ad populum suum conuertetur et singuli ad terram suam fugient 13:15 omnis qui inuentus fuerit occidetur et omnis qui 13:15 Every one that is found shall be thrust through; and

every one that is joined unto them shall fall by the sword.

superuenerit cadet in gladio 13:16 infantes eorum adlident 13:16 Their children also shall be dashed to pieces before

their eyes; their houses shall be spoiled, and their wives ravished.

in oculis eorum diripientur domus eorum et uxores eorum uiolabuntur 13:17 ecce ego suscitabo super eos Medos qui 13:17 Behold, I will stir up the Medes against them, which

shall not regard silver; and as for gold, they shall not delight in it.

argentum non quaerant nec aurum uelint 13:18 sed sagittis

13:18 Their bows also shall dash the young men to pieces; and they shall have no pity on the fruit of the womb; their eyes shall not spare children.

paruulos interficiant et lactantibus uteri non misereantur et super filios non parcat oculus eorum 13:19 et erit Babylon 13:19 And Babylon, the glory of kingdoms, the beauty of

the Chaldees’ excellency, shall be as when God overthrew Sodom and Gomorrah.

illa gloriosa in regnis inclita in superbia Chaldeorum sicut

subuertit Deus Sodomam et Gomorram 13:20 non habitabi- 13:20 It shall never be inhabited, neither shall it be dwelt in from generation to generation: neither shall the Arabian pitch tent there; neither shall the shepherds make their fold

tur usque in finem et non fundabitur usque ad generationem there. et generationem nec ponet ibi tentoria Arabs nec pastores requiescent ibi 13:21 sed requiescent ibi bestiae et reple- 13:21 But wild beasts of the desert shall lie there; and their

houses shall be full of doleful creatures; and owls shall dwell there, and satyrs shall dance there.

buntur domus eorum draconibus et habitabunt ibi strutiones et pilosi saltabunt ibi 13:22 et respondebunt ibi ululae in 13:22 And the wild beasts of the islands shall cry in their

desolate houses, and dragons in their pleasant palaces: and her time is near to come, and her days shall not be pro-

aedibus eius et sirenae in delubris uoluptatis longed.

14:1 prope est ut ueniat tempus eius et dies eius non elon- 14:1 For the LORD will have mercy on Jacob, and will

yet choose Israel, and set them in their own land: and the strangers shall be joined with them, and they shall cleave

gabuntur miserebitur enim Dominus Iacob et eliget adhuc to the house of Jacob. de Israhel et requiescere eos faciet super humum suam adiungetur aduena ad eos et adherebit domui Iacob 14:2 et 14:2 And the people shall take them, and bring them to

their place: and the house of Israel shall possess them in the land of the LORD for servants and handmaids: and they shall take them captives, whose captives they were; and they shall rule over their oppressors.

Biblia Sacra Vulgata

tenebunt eos populi et adducent eos in locum suum et possidebit eos domus Israhel super terram Domini in seruos et ancillas et erunt capientes eos qui se ceperant et subicient

14:3 And it shall come to pass in the day that the LORD exactores suos 14:3 et erit in die illa cum requiem dederit

shall give thee rest from thy sorrow, and from thy fear, and from the hard bondage wherein thou wast made to serve,

tibi Deus a labore tuo et a concussione tua et a seruitute dura

14:4 That thou shalt take up this proverb against the king qua ante seruisti 14:4 sumes parabolam istam contra regem

of Babylon, and say, How hath the oppressor ceased! the golden city ceased!

Babylonis et dices quomodo cessauit exactor quieuit tribu-

14:5 The LORD hath broken the staff of the wicked, and tum 14:5 contriuit Dominus baculum impiorum uirgam do-

the sceptre of the rulers.

14:6 He who smote the people in wrath with a continual minantium 14:6 caedentem populos in indignatione plaga

stroke, he that ruled the nations in anger, is persecuted, and none hindereth.

insanabili subicientem in furore gentes persequentem crude-

14:7 The whole earth is at rest, and is quiet: they break liter 14:7 conquieuit et siluit omnis terra gauisa est et ex-

forth into singing.

14:8 Yea, the fir trees rejoice at thee, and the cedars of ultauit 14:8 abietes quoque laetatae sunt super te et cedri

Lebanon, saying, Since thou art laid down, no feller is come up against us.

Libani ex quo dormisti non ascendit qui succidat nos 14:9

14:9 Hell from beneath is moved for thee to meet thee at thy coming: it stirreth up the dead for thee, even all the chief ones of the earth; it hath raised up from their thrones all the kings of the nations.

infernus subter conturbatus est in occursum aduentus tui suscitauit tibi gigantas omnes principes terrae surrexerunt de so-

14:10 All they shall speak and say unto thee, Art thou also liis suis omnes principes nationum 14:10 uniuersi respon-

become weak as we? art thou become like unto us?

debunt et dicent tibi et tu uulneratus es sicut nos nostri si-

14:11 Thy pomp is brought down to the grave, and the milis effectus es 14:11 detracta est ad inferos superbia tua

noise of thy viols: the worm is spread under thee, and the worms cover thee.

concidit cadauer tuum subter te sternetur tinea et operimen-

14:12 How art thou fallen from heaven, O Lucifer, son of tum tuum erunt uermes 14:12 quomodo cecidisti de caelo the morning! how art thou cut down to the ground, which didst weaken the nations!

lucifer qui mane oriebaris corruisti in terram qui uulnerabas

14:13 For thou hast said in thine heart, I will ascend into gentes 14:13 qui dicebas in corde tuo in caelum conscenheaven, I will exalt my throne above the stars of God: I will sit also upon the mount of the congregation, in the

sides of the north: dam super astra Dei exaltabo solium meum sedebo in monte 14:14 I will ascend above the heights of the clouds; I will testamenti in lateribus aquilonis 14:14 ascendam super al-

be like the most High.

14:15 Yet thou shalt be brought down to hell, to the sides titudinem nubium ero similis Altissimo 14:15 uerumtamen

of the pit.

14:16 They that see thee shall narrowly look upon thee, ad infernum detraheris in profundum laci 14:16 qui te ui-

and consider thee, saying, Is this the man that made the earth to tremble, that did shake kingdoms;

derint ad te inclinabuntur teque prospicient numquid iste est

14:17 That made the world as a wilderness, and destroyed uir qui conturbauit terram qui concussit regna 14:17 qui po-

the cities thereof; that opened not the house of his prisoners?

suit orbem desertum et urbes eius destruxit uinctis eius non

14:18 All the kings of the nations, even all of them, lie in aperuit carcerem 14:18 omnes reges gentium uniuersi dor-

glory, every one in his own house. 14:19 But thou art cast out of thy grave like an abominable mierunt in gloria uir in domo sua 14:19 tu autem proiectus branch, and as the raiment of those that are slain, thrust through with a sword, that go down to the stones of the

pit; as a carcase trodden under feet. es de sepulchro tuo quasi stirps inutilis pollutus et obuolutus qui interfecti sunt gladio et descenderunt ad fundamenta

Prophetia Isaiae

laci quasi cadauer putridum 14:20 non habebis consortium 14:20 Thou shalt not be joined with them in burial, be-

cause thou hast destroyed thy land, and slain thy people: the seed of evildoers shall never be renowned.

neque cum eis in sepultura tu enim terram disperdisti tu populum occidisti non uocabitur in aeternum semen pessimorum 14:21 praeparate filios eius occisioni in iniquitate pa- 14:21 Prepare slaughter for his children for the iniquity of

their fathers; that they do not rise, nor possess the land, nor fill the face of the world with cities.

trum eorum non consurgent nec hereditabunt terram neque implebunt faciem orbis ciuitatum 14:22 et consurgam super 14:22 For I will rise up against them, saith the LORD of

hosts, and cut off from Babylon the name, and remnant, and son, and nephew, saith the LORD.

eos dicit Dominus exercituum et perdam Babylonis nomen

et reliquias et germen et progeniem ait Dominus 14:23 et 14:23 I will also make it a possession for the bittern, and pools of water: and I will sweep it with the besom of destruction, saith the LORD of hosts.

ponam eam in possessionem ericii et in paludes aquarum et scopabo eam in scopa terens dicit Dominus exercituum 14:24 iurauit Dominus exercituum dicens si non ut putaui ita 14:24 The LORD of hosts hath sworn, saying, Surely as

I have thought, so shall it come to pass; and as I have purposed, so shall it stand:

erit et quomodo mente tractaui 14:25 sic eueniet ut conte-

14:25 That I will break the Assyrian in my land, and upon my mountains tread him under foot: then shall his yoke depart from off them, and his burden depart from off their shoulders.

ram Assyrium in terra mea et in montibus meis conculcem eum et auferetur ab eis iugum eius et onus illius ab umero eorum tolletur 14:26 hoc consilium quod cogitaui super om- 14:26 This is the purpose that is purposed upon the whole

earth: and this is the hand that is stretched out upon all the nations.

nem terram et haec est manus extenta super uniuersas gentes 14:27 Dominus enim exercituum decreuit et quis poterit in- 14:27 For the LORD of hosts hath purposed, and who shall

disannul it? and his hand is stretched out, and who shall turn it back?

firmare et manus eius extenta et quis auertet eam 14:28 in

14:28 In the year that king Ahaz died was this burden.

anno quo mortuus est rex Ahaz factum est onus istud 14:29 14:29 Rejoice not thou, whole Palestina, because the rod of him that smote thee is broken: for out of the serpent’s root shall come forth a cockatrice, and his fruit shall be a

ne laeteris Philisthea omnis tu quoniam comminuta est uirga fiery flying serpent. percussoris tui de radice enim colubri egredietur regulus et semen eius absorbens uolucrem 14:30 et pascentur primo- 14:30 And the firstborn of the poor shall feed, and the

needy shall lie down in safety: and I will kill thy root with famine, and he shall slay thy remnant.

geniti pauperum et pauperes fiducialiter requiescent et interire faciam in fame radicem tuam et reliquias tuas interficiam 14:31 ulula porta clama ciuitas prostrata est Philisthea omnis 14:31 Howl, O gate; cry, O city; thou, whole Palestina, art

dissolved: for there shall come from the north a smoke, and none shall be alone in his appointed times.

ab aquilone enim fumus uenit et non est qui effugiat agmen eius 14:32 et quid respondebitur nuntiis gentis quia Domi- 14:32 What shall one then answer the messengers of the

nation? That the LORD hath founded Zion, and the poor of his people shall trust in it.

nus fundauit Sion et in ipsa sperabunt pauperes populi eius

15:1 onus Moab quia nocte uastata est Ar Moab conticuit 15:1 The burden of Moab. Because in the night Ar of

Moab is laid waste, and brought to silence; because in the night Kir of Moab is laid waste, and brought to silence;

quia nocte uastatus est murus Moab conticuit 15:2 ascendit

15:2 He is gone up to Bajith, and to Dibon, the high places, to weep: Moab shall howl over Nebo, and over Medeba: on all their heads shall be baldness, and every beard cut off.

domus et Dibon ad excelsa in planctum super Nabo et super Medaba Moab ululabit in cunctis capitibus eius caluitium omnis barba radetur 15:3 in triuiis eius accincti sunt sacco 15:3 In their streets they shall gird themselves with sack-

cloth: on the tops of their houses, and in their streets, every one shall howl, weeping abundantly.

Biblia Sacra Vulgata

super tecta eius et in plateis eius omnis ululat descendit in

15:4 And Heshbon shall cry, and Elealeh: their voice shall fletum 15:4 clamauit Esebon et Eleale usque Iasa audita est be heard even unto Jahaz: therefore the armed soldiers of Moab shall cry out; his life shall be grievous unto him.

uox eorum super hoc expediti Moab ululabunt anima eius

15:5 My heart shall cry out for Moab; his fugitives shall ululabit sibi 15:5 cor meum ad Moab clamabit uectes eius flee unto Zoar, an heifer of three years old: for by the mounting up of Luhith with weeping shall they go it up; for in the way of Horonaim they shall raise up a cry of destruction.

usque ad Segor uitulam conternantem per ascensum enim Luith flens ascendet et in uia Oronaim clamorem contritio-

15:6 For the waters of Nimrim shall be desolate: for the nis leuabunt 15:6 aquae enim Nemrim desertae erunt quia

hay is withered away, the grass faileth, there is no green thing.

aruit herba defecit germen uiror omnis interiit 15:7 secun-

15:7 Therefore the abundance they have gotten, and that which they have laid up, shall they carry away to the brook of the willows.

dum magnitudinem operis et uisitatio eorum ad torrentem

15:8 For the cry is gone round about the borders of Moab; salicum ducent eos 15:8 quoniam circumiit clamor termi-

the howling thereof unto Eglaim, and the howling thereof unto Beerelim.

num Moab usque ad Gallim ululatus eius et usque ad puteum

15:9 For the waters of Dimon shall be full of blood: for I Helim clamor eius 15:9 quia aquae Dibon repletae sunt san-

will bring more upon Dimon, lions upon him that escapeth of Moab, and upon the remnant of the land.

guine ponam enim super Dibon additamenta his qui fugerint de Moab leonem et reliquiis terrae

16:1 Send ye the lamb to the ruler of the land from Sela to 16:1 emitte agnum dominatorem terrae de Petra deserti ad

the wilderness, unto the mount of the daughter of Zion.

16:2 For it shall be, that, as a wandering bird cast out of montem filiae Sion 16:2 et erit sicut auis fugiens et pulli

the nest, so the daughters of Moab shall be at the fords of Arnon.

de nido auolantes sic erunt filiae Moab in transcensu Arnon

16:3 Take counsel, execute judgment; make thy shadow 16:3 ini consilium coge concilium pone quasi noctem umb-

as the night in the midst of the noonday; hide the outcasts; bewray not him that wandereth.

ram tuam in meridie absconde fugientes et uagos ne prodas

16:4 Let mine outcasts dwell with thee, Moab; be thou a 16:4 habitabunt apud te profugi mei Moab esto latibulum eo-

covert to them from the face of the spoiler: for the extortioner is at an end, the spoiler ceaseth, the oppressors are

consumed out of the land. rum a facie uastatoris finitus est enim puluis consummatus 16:5 And in mercy shall the throne be established: and he est miser defecit qui conculcabat terram 16:5 et praepara-

shall sit upon it in truth in the tabernacle of David, judging, and seeking judgment, and hasting righteousness.

bitur in misericordia solium et sedebit super eum in ueritate in tabernaculo Dauid iudicans et quaerens iudicium et ue-

16:6 We have heard of the pride of Moab; he is very proud: lociter reddens quod iustum est 16:6 audiuimus superbiam

even of his haughtiness, and his pride, and his wrath: but his lies shall not be so.

Moab superbus est ualde superbia eius et arrogantia eius et

16:7 Therefore shall Moab howl for Moab, every one shall indignatio eius plus quam fortitudo eius 16:7 idcirco ulula-

howl: for the foundations of Kirhareseth shall ye mourn; surely they are stricken.

bit Moab ad Moab uniuersus ululabit his qui laetantur super

16:8 For the fields of Heshbon languish, and the vine of muro cocti lateris loquimini plagas suas 16:8 quoniam su-

Sibmah: the lords of the heathen have broken down the principal plants thereof, they are come even unto Jazer, they wandered through the wilderness: her branches are stretched out, they are gone over the sea.

burbana Esebon deserta sunt et uinea Sabama domini gentium exciderunt flagella eius usque ad Iazer peruenerunt errauerunt in deserto propagines eius relictae sunt transierunt

16:9 Therefore I will bewail with the weeping of Jazer the mare 16:9 super hoc plorabo in fletu Iazer uineam Sabama vine of Sibmah: I will water thee with my tears, O Heshbon, and Elealeh: for the shouting for thy summer fruits and for thy harvest is fallen.

Prophetia Isaiae

inebriabo te lacrima mea Esebon et Eleale quoniam super uindemiam tuam et super messem tuam uox calcantium inruit 16:10 et auferetur laetitia et exultatio de Carmelo et in 16:10 And gladness is taken away, and joy out of the plen-

tiful field; and in the vineyards there shall be no singing, neither shall there be shouting: the treaders shall tread out no wine in their presses; I have made their vintage shouting to cease.

uineis non exultabit neque iubilabit uinum in torculari non calcabit qui calcare consueuerat uocem calcantium abstuli 16:11 super hoc uenter meus ad Moab quasi cithara sonabit 16:11 Wherefore my bowels shall sound like an harp for

Moab, and mine inward parts for Kirharesh.

et uiscera mea ad murum cocti lateris 16:12 et erit cum ap- 16:12 And it shall come to pass, when it is seen that Moab

is weary on the high place, that he shall come to his sanctuary to pray; but he shall not prevail.

paruerit quod laborauit Moab super excelsis suis ingredietur ad sancta sua ut obsecret et non ualebit 16:13 hoc uerbum 16:13 This is the word that the LORD hath spoken con-

cerning Moab since that time.

quod locutus est Dominus ad Moab ex tunc 16:14 et nunc 16:14 But now the LORD hath spoken, saying, Within

three years, as the years of an hireling, and the glory of Moab shall be contemned, with all that great multitude;

locutus est Dominus dicens in tribus annis quasi anni mer- and the remnant shall be very small and feeble. cennarii auferetur gloria Moab super omni populo multo et relinquetur paruus et modicus nequaquam multus

17:1 onus Damasci ecce Damascus desinet esse ciuitas et 17:1 The burden of Damascus. Behold, Damascus is taken

away from being a city, and it shall be a ruinous heap.

erit sicut aceruus lapidum in ruina 17:2 derelictae ciuitates 17:2 The cities of Aroer are forsaken: they shall be for

flocks, which shall lie down, and none shall make them afraid.

Aroer gregibus erunt et requiescent ibi et non erit qui exterreat 17:3 et cessabit adiutorium ab Ephraim et regnum 17:3 The fortress also shall cease from Ephraim, and the

kingdom from Damascus, and the remnant of Syria: they shall be as the glory of the children of Israel, saith the

a Damasco et reliquiae Syriae sicut gloria filiorum Israhel LORD of hosts. erunt dicit Dominus exercituum 17:4 et erit in die illa adte- 17:4 And in that day it shall come to pass, that the glory of

Jacob shall be made thin, and the fatness of his flesh shall wax lean.

nuabitur gloria Iacob et pingue carnis eius marcescet 17:5

17:5 And it shall be as when the harvestman gathereth the corn, and reapeth the ears with his arm; and it shall be as he that gathereth ears in the valley of Rephaim.

et erit sicut congregans in messe quod restiterit et brachium eius spicas leget et erit sicut quaerens spicas in ualle Rafaim 17:6 et relinquetur in eo sicut racemus et sicut excussio oleae 17:6 Yet gleaning grapes shall be left in it, as the shaking

of an olive tree, two or three berries in the top of the uppermost bough, four or five in the outmost fruitful branches

duarum aut trium oliuarum in summitate rami siue quattuor thereof, saith the LORD God of Israel. aut quinque in cacuminibus eius fructus eius dicit Dominus Deus Israhel 17:7 in die illa inclinabitur homo ad factorem 17:7 At that day shall a man look to his Maker, and his

eyes shall have respect to the Holy One of Israel.

suum et oculi eius ad Sanctum Israhel respicient 17:8 et non 17:8 And he shall not look to the altars, the work of his

hands, neither shall respect that which his fingers have made, either the groves, or the images.

inclinabitur ad altaria quae fecerunt manus eius et quae operati sunt digiti eius non respiciet lucos et delubra 17:9 in 17:9 In that day shall his strong cities be as a forsaken

bough, and an uppermost branch, which they left because of the children of Israel: and there shall be desolation.

die illa erunt ciuitates fortitudinis eius derelictae sicut aratra et segetes quae derelictae sunt a facie filiorum Israhel et erit deserta 17:10 quia oblita es Dei saluatoris tui et Fortis ad- 17:10 Because thou hast forgotten the God of thy salva-

tion, and hast not been mindful of the rock of thy strength, therefore shalt thou plant pleasant plants, and shalt set it

iutoris tui non es recordata propterea plantabis plantationem with strange slips:

Biblia Sacra Vulgata

17:11 In the day shalt thou make thy plant to grow, and in fidelem et germen alienum seminabis 17:11 in die planta-

the morning shalt thou make thy seed to flourish: but the harvest shall be a heap in the day of grief and of desperate

sorrow. tionis tuae labrusca et mane semen tuum florebit ablata est 17:12 Woe to the multitude of many people, which make messis in die hereditatis et dolebit grauiter 17:12 uae mul-

a noise like the noise of the seas; and to the rushing of nations, that make a rushing like the rushing of mighty

waters! titudo populorum multorum ut multitudo maris sonantis et 17:13 The nations shall rush like the rushing of many wa- tumultus turbarum sicut sonitus aquarum multarum 17:13

ters: but God shall rebuke them, and they shall flee far off, and shall be chased as the chaff of the mountains before

the wind, and like a rolling thing before the whirlwind. sonabunt populi sicut sonitus aquarum inundantium et incre-

pabit eum et fugiet procul et rapietur sicut puluis montium a

17:14 And behold at eveningtide trouble; and before the facie uenti et sicut turbo coram tempestate 17:14 in tempore

morning he is not. This is the portion of them that spoil us, and the lot of them that rob us.

uespere et ecce turbatio in matutino et non subsistet haec est pars eorum qui uastauerunt nos et sors diripientium nos

18:1 Woe to the land shadowing with wings, which is be- 18:1 uae terrae cymbalo alarum quae est trans flumina Ae-

yond the rivers of Ethiopia:

18:2 That sendeth ambassadors by the sea, even in vessels thiopiae 18:2 qui mittit in mari legatos et in uasis papyri

of bulrushes upon the waters, saying, Go, ye swift messengers, to a nation scattered and peeled, to a people terrible from their beginning hitherto; a nation meted out and trodden down, whose land the rivers have spoiled!

super aquas ite angeli ueloces ad gentem conuulsam et dilaceratam ad populum terribilem post quem non est alius gentem expectantem expectantem et conculcatam cuius diripue-

18:3 All ye inhabitants of the world, and dwellers on the runt flumina terram eius 18:3 omnes habitatores orbis qui

earth, see ye, when he lifteth up an ensign on the mountains; and when he bloweth a trumpet, hear ye.

moramini in terra cum eleuatum fuerit signum in montibus

18:4 For so the LORD said unto me, I will take my rest, uidebitis et clangorem tubae audietis 18:4 quia haec dicit

and I will consider in my dwelling place like a clear heat upon herbs, and like a cloud of dew in the heat of harvest.

Dominus ad me quiescam et considerabo in loco meo sicut meridiana lux clara est et sicut nubes roris in die messis

18:5 For afore the harvest, when the bud is perfect, and 18:5 ante messem enim totus effloruit et inmatura perfectio

the sour grape is ripening in the flower, he shall both cut off the sprigs with pruning hooks, and take away and cut

down the branches. germinabit et praecidentur ramusculi eius falcibus et quae 18:6 They shall be left together unto the fowls of the derelicta fuerint abscidentur excutientur 18:6 et relinquen-

mountains, and to the beasts of the earth: and the fowls shall summer upon them, and all the beasts of the earth

shall winter upon them. tur simul auibus montium et bestiis terrae et aestate perpetua

erunt super eum uolucres et omnes bestiae terrae super illum

18:7 In that time shall the present be brought unto the hiemabunt 18:7 in tempore illo deferetur munus Domino

LORD of hosts of a people scattered and peeled, and from a people terrible from their beginning hitherto; a nation meted out and trodden under foot, whose land the rivers have spoiled, to the place of the name of the LORD of hosts, the mount Zion.

exercituum a populo diuulso et dilacerato a populo terribili post quem non fuit alius a gente expectante expectante et conculcata cuius diripuerunt flumina terram eius ad locum nominis Domini exercituum montem Sion

19:1 The burden of Egypt. Behold, the LORD rideth upon 19:1 onus Aegypti ecce Dominus ascendet super nubem a swift cloud, and shall come into Egypt: and the idols of Egypt shall be moved at his presence, and the heart of

Egypt shall melt in the midst of it. leuem et ingredietur Aegyptum et mouebuntur simulacra Aegypti a facie eius et cor Aegypti tabescet in medio eius

Prophetia Isaiae

19:2 et concurrere faciam Aegyptios aduersum Aegyptios et 19:2 And I will set the Egyptians against the Egyptians:

and they shall fight every one against his brother, and every one against his neighbour; city against city, and king-

pugnabit uir contra fratrem suum et uir contra amicum suum dom against kingdom.

ciuitas aduersus ciuitatem regnum aduersus regnum 19:3 19:3 And the spirit of Egypt shall fail in the midst thereof; and I will destroy the counsel thereof: and they shall seek to the idols, and to the charmers, and to them that have

et disrumpetur spiritus Aegypti in uisceribus eius et consi- familiar spirits, and to the wizards. lium eius praecipitabo et interrogabunt simulacra sua et diuinos suos et pythones et ariolos 19:4 et tradam Aegyptum 19:4 And the Egyptians will I give over into the hand of a

cruel lord; and a fierce king shall rule over them, saith the Lord, the LORD of hosts.

in manu dominorum crudelium et rex fortis dominabitur eo-

rum ait Dominus Deus exercituum 19:5 et arescet aqua de 19:5 And the waters shall fail from the sea, and the river

shall be wasted and dried up.

mari et fluuius desolabitur atque siccabitur 19:6 et deficient 19:6 And they shall turn the rivers far away; and the

brooks of defence shall be emptied and dried up: the reeds and flags shall wither.

flumina adtenuabuntur et siccabuntur riui aggerum calamus et iuncus marcescet 19:7 nudabitur alueus riui a fonte suo 19:7 The paper reeds by the brooks, by the mouth of the

brooks, and every thing sown by the brooks, shall wither, be driven away, and be no more.

et omnis sementis inrigua siccabitur arescet et non erit 19:8

19:8 The fishers also shall mourn, and all they that cast angle into the brooks shall lament, and they that spread nets upon the waters shall languish.

et maerebunt piscatores et lugebunt omnes mittentes in flumen hamum et expandentes rete super faciem aquae marcescent 19:9 confundentur qui operabantur linum pectentes et 19:9 Moreover they that work in fine flax, and they that

weave networks, shall be confounded.

texentes subtilia 19:10 et erunt inrigua eius flaccentia om- 19:10 And they shall be broken in the purposes thereof, all

that make sluices and ponds for fish.

nes qui faciebant lacunas ad capiendos pisces 19:11 stulti 19:11 Surely the princes of Zoan are fools, the counsel of

the wise counsellors of Pharaoh is become brutish: how say ye unto Pharaoh, I am the son of the wise, the son of

principes Taneos sapientes consiliarii Pharao dederunt con- ancient kings? silium insipiens quomodo dicetis Pharaoni filius sapientium ego filius regum antiquorum 19:12 ubi sunt nunc sapientes 19:12 Where are they? where are thy wise men? and let

them tell thee now, and let them know what the LORD of hosts hath purposed upon Egypt.

tui adnuntient tibi et indicent quid cogitauerit Dominus exercituum super Aegyptum 19:13 stulti facti sunt principes 19:13 The princes of Zoan are become fools, the princes

of Noph are deceived; they have also seduced Egypt, even they that are the stay of the tribes thereof.

Taneos emarcuerunt principes Mempheos deceperunt Aegyptum angulum populorum eius 19:14 Dominus miscuit in 19:14 The LORD hath mingled a perverse spirit in the

midst thereof: and they have caused Egypt to err in every work thereof, as a drunken man staggereth in his vomit.

medio eius spiritum uertiginis et errare fecerunt Aegyptum in omni opere suo sicut errat ebrius et uomens 19:15 et non 19:15 Neither shall there be any work for Egypt, which the

head or tail, branch or rush, may do.

erit Aegypto opus quod faciat caput et caudam incuruantem et refrenantem 19:16 in die illa erit Aegyptus quasi mulieres 19:16 In that day shall Egypt be like unto women: and it

shall be afraid and fear because of the shaking of the hand of the LORD of hosts, which he shaketh over it.

et stupebunt et timebunt a facie commotionis manus Domini exercituum quam ipse mouebit super eam 19:17 et erit terra 19:17 And the land of Judah shall be a terror unto Egypt,

every one that maketh mention thereof shall be afraid in himself, because of the counsel of the LORD of hosts,

Iuda Aegypto in festiuitatem omnis qui illius fuerit recor- which he hath determined against it. datus pauebit a facie consilii Domini exercituum quod ipse cogitauit super eam 19:18 in die illa erunt quinque ciuita- 19:18 In that day shall five cities in the land of Egypt speak

the language of Canaan, and swear to the LORD of hosts; one shall be called, The city of destruction.

Biblia Sacra Vulgata

tes in terra Aegypti loquentes lingua Chanaan et iurantes per

19:19 In that day shall there be an altar to the LORD in Dominum exercituum ciuitas Solis uocabitur una 19:19 in

the midst of the land of Egypt, and a pillar at the border thereof to the LORD.

die illa erit altare Domini in medio terrae Aegypti et titulus

19:20 And it shall be for a sign and for a witness unto the iuxta terminum eius Domini 19:20 et erit in signum et in te-

LORD of hosts in the land of Egypt: for they shall cry unto the LORD because of the oppressors, and he shall send them a saviour, and a great one, and he shall deliver them.

stimonium Domino exercituum in terra Aegypti clamabunt enim ad Dominum a facie tribulantis et mittet eis saluatorem

19:21 And the LORD shall be known to Egypt, and the et propugnatorem qui liberet eos 19:21 et cognoscetur Do-

Egyptians shall know the LORD in that day, and shall do sacrifice and oblation; yea, they shall vow a vow unto the

LORD, and perform it. minus ab Aegypto et cognoscent Aegyptii Dominum in die

illa et colent eum in hostiis et muneribus et uota uouebunt

19:22 And the LORD shall smite Egypt: he shall smite Domino et soluent 19:22 et percutiet Dominus Aegyptum

and heal it: and they shall return even to the LORD, and he shall be intreated of them, and shall heal them.

plaga et sanabit eam et reuertentur ad Dominum et placabi-

19:23 In that day shall there be a highway out of Egypt to tur eis et sanabit eos 19:23 in die illa erit uia de Aegypto

Assyria, and the Assyrian shall come into Egypt, and the Egyptian into Assyria, and the Egyptians shall serve with

the Assyrians. in Assyrios et intrabit Assyrius Aegyptum et Aegyptius in 19:24 In that day shall Israel be the third with Egypt and Assyrios et seruient Aegyptii Assur 19:24 in die illa erit Is-

with Assyria, even a blessing in the midst of the land:

rahel tertius Aegyptio et Assyrio benedictio in medio terrae

19:25 Whom the LORD of hosts shall bless, saying, 19:25 cui benedixit Dominus exercituum dicens benedictus

Blessed be Egypt my people, and Assyria the work of my hands, and Israel mine inheritance.

populus meus Aegypti et opus manuum mearum Assyrio hereditas autem mea Israhel

20:1 In the year that Tartan came unto Ashdod, (when Sar- 20:1 in anno quo ingressus est Tharthan in Azotum cum gon the king of Assyria sent him,) and fought against Ashdod, and took it;

misisset eum Sargon rex Assyriorum et pugnasset contra

20:2 At the same time spake the LORD by Isaiah the son Azotum et cepisset eam 20:2 in tempore illo locutus est Do-

of Amoz, saying, Go and loose the sackcloth from off thy loins, and put off thy shoe from thy foot. And he did so,

walking naked and barefoot. minus in manu Isaiae filii Amos dicens uade et solue saccum

de lumbis tuis et calciamenta tua tolle de pedibus tuis et fecit

20:3 And the LORD said, Like as my servant Isaiah hath sic uadens nudus et disculciatus 20:3 et dixit Dominus sic-

walked naked and barefoot three years for a sign and wonder upon Egypt and upon Ethiopia;

ut ambulauit seruus meus Isaias nudus et disculciatus trium annorum signum et portentum erit super Aegyptum et su-

20:4 So shall the king of Assyria lead away the Egyp- per Aethiopiam 20:4 sic minabit rex Assyriorum captiuita-

tians prisoners, and the Ethiopians captives, young and old, naked and barefoot, even with their buttocks uncov-

ered, to the shame of Egypt. tem Aegypti et transmigrationem Aethiopiae iuuenum et se-

num nudam et disculciatam discopertis natibus ignominiam

20:5 And they shall be afraid and ashamed of Ethiopia Aegypti 20:5 et timebunt et confundentur ab Aethiopia spe

their expectation, and of Egypt their glory.

20:6 And the inhabitant of this isle shall say in that day, sua et ab Aegypto gloria sua 20:6 et dicet habitator insulae

Behold, such is our expectation, whither we flee for help to be delivered from the king of Assyria: and how shall we

escape? huius in die illa ecce haec erat spes nostra ad quos confugimus in auxilium ut liberaret nos a facie regis Assyriorum et

Prophetia Isaiae

quomodo effugere poterimus nos

21:1 onus deserti maris sicut turbines ab africo ueniunt 21:1 The burden of the desert of the sea. As whirlwinds in

the south pass through; so it cometh from the desert, from a terrible land.

de deserto uenit de terra horribili 21:2 uisio dura nuntiata

21:2 A grievous vision is declared unto me; the treacherous dealer dealeth treacherously, and the spoiler spoileth. Go up, O Elam: besiege, O Media; all the sighing thereof have I made to cease.

est mihi qui incredulus est infideliter agit et qui depopulator est uastat ascende Aelam obside Mede omnem gemitum eius cessare feci 21:3 propterea repleti sunt lumbi mei 21:3 Therefore are my loins filled with pain: pangs have

taken hold upon me, as the pangs of a woman that travaileth: I was bowed down at the hearing of it; I was dis-

dolore angustia possedit me sicut angustia parientis corrui mayed at the seeing of it.

cum audirem conturbatus sum cum uiderem 21:4 emarcuit 21:4 My heart panted, fearfulness affrighted me: the night

of my pleasure hath he turned into fear unto me.

cor meum tenebrae stupefecerunt me Babylon dilecta mea

posita est mihi in miraculum 21:5 pone mensam contemp- 21:5 Prepare the table, watch in the watchtower, eat, drink:

arise, ye princes, and anoint the shield.

lare in specula comedentes bibentes surgite principes arripite clypeum 21:6 haec enim dixit mihi Dominus uade et pone 21:6 For thus hath the LORD said unto me, Go, set a

watchman, let him declare what he seeth.

speculatorem et quodcumque uiderit adnuntiet 21:7 et ui- 21:7 And he saw a chariot with a couple of horsemen, a

chariot of asses, and a chariot of camels; and he hearkened diligently with much heed:

dit currum duorum equitum ascensorem asini et ascensorem cameli et contemplatus est diligenter multo intuitu 21:8 et 21:8 And he cried, A lion: My lord, I stand continually

upon the watchtower in the daytime, and I am set in my ward whole nights:

clamauit leo super specula Domini ego sum stans iugiter per diem et super custodiam meam ego sum stans totis noctibus 21:9 ecce iste uenit ascensor uir bigae equitum et respondit 21:9 And, behold, here cometh a chariot of men, with a

couple of horsemen. And he answered and said, Babylon is fallen, is fallen; and all the graven images of her gods

et dixit cecidit cecidit Babylon et omnia sculptilia deorum he hath broken unto the ground. eius contrita sunt in terram 21:10 tritura mea et fili areae 21:10 O my threshing, and the corn of my floor: that which

I have heard of the LORD of hosts, the God of Israel, have I declared unto you.

meae quae audiui a Domino exercituum Deo Israhel adnuntiaui uobis 21:11 onus Duma ad me clamat ex Seir custos 21:11 The burden of Dumah. He calleth to me out of Seir,

Watchman, what of the night? Watchman, what of the night?

quid de nocte custos quid de nocte 21:12 dixit custos uenit

21:12 The watchman said, The morning cometh, and also the night: if ye will enquire, enquire ye: return, come.

mane et nox si quaeritis quaerite conuertimini uenite 21:13 _{21:13 The burden upon Arabia. In the forest in Arabia}

shall ye lodge, O ye travelling companies of Dedanim.

onus in Arabia in saltu ad uesperam dormietis in semitis Dodanim 21:14 occurrentes sitienti ferte aquam qui habitatis 21:14 The inhabitants of the land of Tema brought water

to him that was thirsty, they prevented with their bread him that fled.

terram austri cum panibus occurrite fugienti 21:15 a facie

21:15 For they fled from the swords, from the drawn sword, and from the bent bow, and from the grievousness of war.

enim gladiorum fugerunt a facie gladii inminentis a facie arcus extenti a facie grauis proelii 21:16 quoniam haec dicit 21:16 For thus hath the LORD said unto me, Within a year,

according to the years of an hireling, and all the glory of Kedar shall fail:

Dominus ad me adhuc in uno anno quasi in anno mercennarii et auferetur omnis gloria Cedar 21:17 et reliquiae numeri 21:17 And the residue of the number of archers, the

mighty men of the children of Kedar, shall be diminished: for the LORD God of Israel hath spoken it.

sagittariorum fortium de filiis Cedar inminuentur Dominus enim Deus Israhel locutus est

Biblia Sacra Vulgata

22:1 The burden of the valley of vision. What aileth thee 22:1 onus uallis Visionis quidnam tibi quoque est quia as-

now, that thou art wholly gone up to the housetops? 22:2 Thou that art full of stirs, a tumultuous city, joyous cendisti et tu omnis in tecta 22:2 clamoris plena urbs frecity: thy slain men are not slain with the sword, nor dead in battle.

quens ciuitas exultans interfecti tui non interfecti gladio nec

22:3 All thy rulers are fled together, they are bound by mortui in bello 22:3 cuncti principes tui fugerunt simul du-

the archers: all that are found in thee are bound together, which have fled from far.

reque ligati sunt omnes qui inuenti sunt uincti sunt pariter

22:4 Therefore said I, Look away from me; I will weep procul fugerunt 22:4 propterea dixi recedite a me amare

bitterly, labour not to comfort me, because of the spoiling of the daughter of my people.

flebo nolite incumbere ut consolemini me super uastitate fi-

22:5 For it is a day of trouble, and of treading down, and liae populi mei 22:5 dies enim interfectionis et conculca-

of perplexity by the Lord GOD of hosts in the valley of vision, breaking down the walls, and of crying to the moun-

tains. tionis et fletuum Domino Deo exercituum in ualle Visionis 22:6 And Elam bare the quiver with chariots of men and scrutans murum et magnificus super montem 22:6 et Aelam

horsemen, and Kir uncovered the shield.

sumpsit faretram currum hominis equitis et parietem nuda-

22:7 And it shall come to pass, that thy choicest valleys uit clypeus 22:7 et erunt electae ualles tuae plenae quadri-

shall be full of chariots, and the horsemen shall set themselves in array at the gate.

garum et equites ponent sedes suas in porta 22:8 et reue-

22:8 And he discovered the covering of Judah, and thou didst look in that day to the armour of the house of the forest.

labitur operimentum Iudae et uidebis in die illa armamenta-

22:9 Ye have seen also the breaches of the city of David, rium domus saltus 22:9 et scissuras ciuitatis Dauid uidebi-

that they are many: and ye gathered together the waters of the lower pool.

tis quia multiplicatae sunt et congregastis aquas piscinae in-

22:10 And ye have numbered the houses of Jerusalem, and ferioris 22:10 et domos Hierusalem numerastis et destruxi-

the houses have ye broken down to fortify the wall. 22:11 Ye made also a ditch between the two walls for the stis domos ad muniendum murum 22:11 et lacum fecistis water of the old pool: but ye have not looked unto the maker thereof, neither had respect unto him that fashioned

it long ago. inter duos muros et aquam piscinae ueteris et non suspexis-

tis ad eum qui fecerat eam et operatorem eius de longe non

22:12 And in that day did the Lord GOD of hosts call to uidistis 22:12 et uocauit Dominus Deus exercituum in die

weeping, and to mourning, and to baldness, and to girding with sackcloth:

illa ad fletum et ad planctum ad caluitium et ad cingulum

22:13 And behold joy and gladness, slaying oxen, and sacci 22:13 et ecce gaudium et laetitia occidere uitulos et

killing sheep, eating flesh, and drinking wine: let us eat and drink; for to morrow we shall die.

iugulare arietes comedere carnes et bibere uinum comeda-

22:14 And it was revealed in mine ears by the LORD of mus et bibamus cras enim moriemur 22:14 et reuelata est

hosts, Surely this iniquity shall not be purged from you till ye die, saith the Lord GOD of hosts.

in auribus meis Domini exercituum si dimittetur iniquitas haec uobis donec moriamini dicit Dominus Deus exercituum

22:15 Thus saith the Lord GOD of hosts, Go, get thee unto 22:15 haec dicit Dominus Deus exercituum uade ingredere ad

this treasurer, even unto Shebna, which is over the house, and say,

eum qui habitat in tabernaculo ad Sobnam praepositum tem-

22:16 What hast thou here? and whom hast thou here, pli 22:16 quid tu hic aut quasi quis hic quia excidisti tibi

that thou hast hewed thee out a sepulchre here, as he that heweth him out a sepulchre on high, and that graveth an

habitation for himself in a rock? hic sepulchrum excidisti in excelso memoriam diligenter in 22:17 Behold, the LORD will carry thee away with a petra tabernaculum tibi 22:17 ecce Dominus asportari te fa-

mighty captivity, and will surely cover thee.

ciet sicut asportatur gallus gallinacius et quasi amictum sic

Prophetia Isaiae

subleuabit te 22:18 coronans coronabit te tribulatione quasi 22:18 He will surely violently turn and toss thee like a

ball into a large country: there shalt thou die, and there the chariots of thy glory shall be the shame of thy lord’s

pilam mittet te in terram latam et spatiosam ibi morieris et ibi house. erit currus gloriae tuae ignominia domus Domini tui 22:19 22:19 And I will drive thee from thy station, and from thy

state shall he pull thee down.

et expellam te de statione tua et de ministerio tuo deponam te 22:20 et erit in die illa uocabo seruum meum Eliachim filium 22:20 And it shall come to pass in that day, that I will call

my servant Eliakim the son of Hilkiah:

Helciae 22:21 et induam illum tunicam tuam et cingulo tuo 22:21 And I will clothe him with thy robe, and strengthen

him with thy girdle, and I will commit thy government into his hand: and he shall be a father to the inhabitants of

confortabo eum et potestatem tuam dabo in manu eius et erit Jerusalem, and to the house of Judah. quasi pater habitantibus Hierusalem et domui Iuda 22:22 et 22:22 And the key of the house of David will I lay upon

his shoulder; so he shall open, and none shall shut; and he shall shut, and none shall open.

dabo clauem domus Dauid super umerum eius et aperiet et non erit qui claudat et claudet et non erit qui aperiat 22:23 22:23 And I will fasten him as a nail in a sure place; and

he shall be for a glorious throne to his father’s house.

et figam illum paxillum in loco fideli et erit in solium gloriae domui patris sui 22:24 et suspendent super eum omnem 22:24 And they shall hang upon him all the glory of his

father’s house, the offspring and the issue, all vessels of small quantity, from the vessels of cups, even to all the

gloriam domus patris eius uasorum diuersa genera omne uas vessels of flagons. paruulum a uasis craterarum usque ad omne uas musicorum 22:25 in die illo dicit Dominus exercituum auferetur paxillus 22:25 In that day, saith the LORD of hosts, shall the nail

that is fastened in the sure place be removed, and be cut down, and fall; and the burden that was upon it shall be

qui fixus fuerat in loco fideli et frangetur et cadet et peribit cut off: for the LORD hath spoken it. quod pependerat in eo quia Dominus locutus est

23:1 onus Tyri ululate naues maris quia uastata est do- 23:1 The burden of Tyre. Howl, ye ships of Tarshish; for

it is laid waste, so that there is no house, no entering in: from the land of Chittim it is revealed to them.

mus unde uenire consueuerant de terra Cetthim reuelatum est eis 23:2 tacete qui habitatis in insula negotiatio Sidonis 23:2 Be still, ye inhabitants of the isle; thou whom the

merchants of Zidon, that pass over the sea, have replenished.

transfretantes mare repleuerunt te 23:3 in aquis multis se-

23:3 And by great waters the seed of Sihor, the harvest of the river, is her revenue; and she is a mart of nations.

men Nili messis fluminis fruges eius et facta est negotiatio gentium 23:4 erubesce Sidon ait enim mare fortitudo ma- 23:4 Be thou ashamed, O Zidon: for the sea hath spo-

ken, even the strength of the sea, saying, I travail not, nor bring forth children, neither do I nourish up young men,

ris dicens non parturiui et non peperi et non enutriui iuue- nor bring up virgins. nes nec ad incrementum perduxi uirgines 23:5 cum audi- 23:5 As at the report concerning Egypt, so shall they be

sorely pained at the report of Tyre.

tum fuerit in Aegypto dolebunt cum audierint de Tyro 23:6 23:6 Pass ye over to Tarshish; howl, ye inhabitants of the

isle.

transite maria ululate qui habitatis in insula 23:7 numquid 23:7 Is this your joyous city, whose antiquity is of ancient

days? her own feet shall carry her afar off to sojourn.

non haec uestra est quae gloriabatur a diebus pristinis in antiquitate sua ducent eam pedes sui longe ad peregrinan-

dum 23:8 quis cogitauit hoc super Tyrum quondam corona- 23:8 Who hath taken this counsel against Tyre, the crowning city, whose merchants are princes, whose traffickers are the honourable of the earth?

tam cuius negotiatores principes institores eius incliti terrae 23:9 Dominus exercituum cogitauit hoc ut detraheret super- 23:9 The LORD of hosts hath purposed it, to stain the

pride of all glory, and to bring into contempt all the honourable of the earth.

biam omnis gloriae et ad ignominiam deduceret uniuersos

Biblia Sacra Vulgata

23:10 Pass through thy land as a river, O daughter of inclitos terrae 23:10 transi terram tuam quasi flumen filia

Tarshish: there is no more strength. 23:11 He stretched out his hand over the sea, he shook the maris non est cingulum ultra tibi 23:11 manum suam exkingdoms: the LORD hath given a commandment against the merchant city, to destroy the strong holds thereof.

tendit super mare conturbauit regna Dominus mandauit ad-

23:12 And he said, Thou shalt no more rejoice, O thou uersum Chanaan ut contereret fortes eius 23:12 et dixit non

oppressed virgin, daughter of Zidon: arise, pass over to Chittim; there also shalt thou have no rest.

adicies ultra ut glorieris calumniam sustinens uirgo filia Sidonis in Cetthim consurgens transfreta ibi quoque non erit

23:13 Behold the land of the Chaldeans; this people was requies tibi 23:13 ecce terra Chaldeorum talis populus non

not, till the Assyrian founded it for them that dwell in the wilderness: they set up the towers thereof, they raised up

the palaces thereof; and he brought it to ruin. fuit Assur fundauit eam in captiuitatem transduxerunt robu-

stos eius suffoderunt domos eius posuerunt eam in ruinam

23:14 Howl, ye ships of Tarshish: for your strength is laid 23:14 ululate naues maris quia deuastata est fortitudo uestra

waste.

23:15 And it shall come to pass in that day, that Tyre shall 23:15 et erit in die illa in obliuione eris o Tyre septuaginta an-

be forgotten seventy years, according to the days of one king: after the end of seventy years shall Tyre sing as an

harlot. nis sicut dies regis unius post septuaginta autem annos erit 23:16 Take an harp, go about the city, thou harlot that hast Tyro quasi canticum meretricis 23:16 sume citharam circui

been forgotten; make sweet melody, sing many songs, that thou mayest be remembered.

ciuitatem meretrix obliuioni tradita bene cane frequenta can-

23:17 And it shall come to pass after the end of seventy ticum ut memoria tui sit 23:17 et erit post septuaginta annos

years, that the LORD will visit Tyre, and she shall turn to her hire, and shall commit fornication with all the king-

doms of the world upon the face of the earth. uisitabit Dominus Tyrum et reducet eam ad mercedes suas et

rursum fornicabitur cum uniuersis regnis terrae super faciem

23:18 And her merchandise and her hire shall be holiness terrae 23:18 et erunt negotiatio eius et mercedes eius sanc-

to the LORD: it shall not be treasured nor laid up; for her merchandise shall be for them that dwell before the

LORD, to eat sufficiently, and for durable clothing. tificatae Domino non condentur neque reponentur quia his qui habitauerint coram Domino erit negotiatio eius ut manducent in saturitatem et uestiantur usque ad uetustatem

24:1 Behold, the LORD maketh the earth empty, and 24:1 ecce Dominus dissipabit terram et nudabit eam et ad-

maketh it waste, and turneth it upside down, and scattereth abroad the inhabitants thereof.

fliget faciem eius et disperget habitatores eius 24:2 et erit

24:2 And it shall be, as with the people, so with the priest; as with the servant, so with his master; as with the maid, so with her mistress; as with the buyer, so with the seller; as with the lender, so with the borrower; as with the taker of usury, so with the giver of usury to him.

sicut populus sic sacerdos et sicut seruus sic dominus eius sicut ancilla sic domina eius sicut emens sic ille qui uendit sicut fenerator sic is qui mutuum accipit sicut qui repe-

24:3 The land shall be utterly emptied, and utterly spoiled: tit sic qui debet 24:3 dissipatione dissipabitur terra et di-

for the LORD hath spoken this word.

reptione praedabitur Dominus enim locutus est uerbum hoc

24:4 The earth mourneth and fadeth away, the world lan- 24:4 luxit et defluxit terra et infirmata est defluxit orbis in-

guisheth and fadeth away, the haughty people of the earth do languish.

firmata est altitudo populi terrae 24:5 et terra interfecta est

24:5 The earth also is defiled under the inhabitants thereof; because they have transgressed the laws, changed the ordinance, broken the everlasting covenant.

ab habitatoribus suis quia transgressi sunt leges mutauerunt

24:6 Therefore hath the curse devoured the earth, and they ius dissipauerunt foedus sempiternum 24:6 propter hoc ma-

that dwell therein are desolate: therefore the inhabitants of the earth are burned, and few men left.

ledictio uorabit terram et peccabunt habitatores eius ideo-

Prophetia Isaiae

que insanient cultores eius et relinquentur homines pauci 24:7 luxit uindemia infirmata est uitis ingemuerunt omnes 24:7 The new wine mourneth, the vine languisheth, all the

merryhearted do sigh.

qui laetabantur corde 24:8 cessauit gaudium tympanorum 24:8 The mirth of tabrets ceaseth, the noise of them that

rejoice endeth, the joy of the harp ceaseth.

quieuit sonitus laetantium conticuit dulcedo citharae 24:9 24:9 They shall not drink wine with a song; strong drink

shall be bitter to them that drink it.

cum cantico non bibent uinum amara erit potio bibentibus illam 24:10 adtrita est ciuitas uanitatis clausa est omnis do- 24:10 The city of confusion is broken down: every house

is shut up, that no man may come in.

mus nullo introeunte 24:11 clamor erit super uino in pla- 24:11 There is a crying for wine in the streets; all joy is

darkened, the mirth of the land is gone.

teis deserta est omnis laetitia translatum est gaudium terrae 24:12 relicta est in urbe solitudo et calamitas opprimet portas 24:12 In the city is left desolation, and the gate is smitten

with destruction.

24:13 quia haec erunt in medio terrae in medio populorum 24:13 When thus it shall be in the midst of the land among

the people, there shall be as the shaking of an olive tree, and as the gleaning grapes when the vintage is done.

quomodo si paucae oliuae quae remanserunt excutiantur ex olea et racemi cum fuerit finita uindemia 24:14 hii leuabunt 24:14 They shall lift up their voice, they shall sing for the

majesty of the LORD, they shall cry aloud from the sea.

uocem suam atque laudabunt cum glorificatus fuerit Dominus hinnient de mari 24:15 propter hoc in doctrinis glorifi- 24:15 Wherefore glorify ye the LORD in the fires, even

the name of the LORD God of Israel in the isles of the sea.

cate Dominum in insulis maris nomen Domini Dei Israhel 24:16 a finibus terrae laudes audiuimus gloriam iusti et dixi 24:16 From the uttermost part of the earth have we heard

songs, even glory to the righteous. But I said, My leanness, my leanness, woe unto me! the treacherous dealers have dealt treacherously; yea, the treacherous dealers have dealt very treacherously.

secretum meum mihi secretum meum mihi uae mihi praeuaricantes praeuaricati sunt et praeuaricatione transgressorum praeuaricati sunt 24:17 formido et fouea et laqueus super te 24:17 Fear, and the pit, and the snare, are upon thee, O

inhabitant of the earth.

qui habitator es terrae 24:18 et erit qui fugerit a uoce formi- 24:18 And it shall come to pass, that he who fleeth from

the noise of the fear shall fall into the pit; and he that cometh up out of the midst of the pit shall be taken in the snare: for the windows from on high are open, and the foundations of the earth do shake.

dinis cadet in foueam et qui se explicuerit de fouea tenebitur laqueo quia cataractae de excelsis apertae sunt et concutientur fundamenta terrae 24:19 confractione confringetur terra 24:19 The earth is utterly broken down, the earth is clean

dissolved, the earth is moved exceedingly.

contritione conteretur terra commotione commouebitur terra 24:20 agitatione agitabitur terra sicut ebrius et auferetur quasi 24:20 The earth shall reel to and fro like a drunkard, and

shall be removed like a cottage; and the transgression thereof shall be heavy upon it; and it shall fall, and not

tabernaculum unius noctis et grauabit eam iniquitas sua et rise again. corruet et non adiciet ut resurgat 24:21 et erit in die illa 24:21 And it shall come to pass in that day, that the LORD

shall punish the host of the high ones that are on high, and the kings of the earth upon the earth.

uisitabit Dominus super militiam caeli in excelso et super reges terrae qui sunt super terram 24:22 et congregabuntur 24:22 And they shall be gathered together, as prisoners are

gathered in the pit, and shall be shut up in the prison, and after many days shall they be visited.

in congregationem unius fascis in lacum et cludentur ibi in carcerem et post multos dies uisitabuntur 24:23 et erubescet 24:23 Then the moon shall be confounded, and the sun

ashamed, when the LORD of hosts shall reign in mount Zion, and in Jerusalem, and before his ancients gloriously.

luna et confundetur sol cum regnauerit Dominus exercituum in monte Sion et in Hierusalem et in conspectu senum suo-

Biblia Sacra Vulgata

rum fuerit glorificatus

25:1 O Lord, thou art my God; I will exalt thee, I will 25:1 Domine Deus meus es tu exaltabo te confitebor no-

praise thy name; for thou hast done wonderful things; thy counsels of old are faithfulness and truth.

mini tuo quoniam fecisti mirabilia cogitationes antiquas fi-

25:2 For thou hast made of a city an heap; of a defenced deles amen 25:2 quia posuisti ciuitatem in tumulum urbem

city a ruin: a palace of strangers to be no city; it shall never be built.

fortem in ruinam domum alienorum ut non sit ciuitas et in

25:3 Therefore shall the strong people glorify thee, the city sempiternum non aedificetur 25:3 super hoc laudabit te poof the terrible nations shall fear thee.

25:4 For thou hast been a strength to the poor, a strength pulus fortis ciuitas gentium robustarum timebit te 25:4 quia

to the needy in his distress, a refuge from the storm, a shadow from the heat, when the blast of the terrible ones

is as a storm against the wall. factus es fortitudo pauperi fortitudo egeno in tribulatione sua

spes a turbine umbraculum ab aestu spiritus enim robusto-

25:5 Thou shalt bring down the noise of strangers, as the rum quasi turbo inpellens parietem 25:5 sicut aestum in siti

heat in a dry place; even the heat with the shadow of a cloud: the branch of the terrible ones shall be brought low.

tumultum alienorum humiliabis et quasi calore sub nube tor-

25:6 And in this mountain shall the LORD of hosts make rente propaginem fortium marcescere facies 25:6 et faciet

unto all people a feast of fat things, a feast of wines on the lees, of fat things full of marrow, of wines on the lees well

refined. Dominus exercituum omnibus populis in monte hoc conui-

uium pinguium conuiuium uindemiae pinguium medullato-

25:7 And he will destroy in this mountain the face of the rum uindemiae defecatae 25:7 et praecipitabit in monte isto

covering cast over all people, and the vail that is spread over all nations.

faciem uinculi conligati super omnes populos et telam quam

25:8 He will swallow up death in victory; and the Lord orditus est super uniuersas nationes 25:8 praecipitabit mor-

GOD will wipe away tears from off all faces; and the rebuke of his people shall he take away from off all the earth:

for the LORD hath spoken it. tem in sempiternum et auferet Dominus Deus lacrimam ab

omni facie et obprobrium populi sui auferet de uniuersa terra

25:9 And it shall be said in that day, Lo, this is our God; we quia Dominus locutus est 25:9 et dicet in die illa ecce Deus

have waited for him, and he will save us: this is the LORD; we have waited for him, we will be glad and rejoice in his

salvation. noster iste expectauimus eum et saluabit nos iste Dominus

sustinuimus eum exultabimus et laetabimur in salutari eius

25:10 For in this mountain shall the hand of the LORD 25:10 quia requiescet manus Domini in monte isto et tritura-

rest, and Moab shall be trodden down under him, even as straw is trodden down for the dunghill.

bitur Moab sub eo sicuti teruntur paleae in plaustro 25:11 et

25:11 And he shall spread forth his hands in the midst of them, as he that swimmeth spreadeth forth his hands to swim: and he shall bring down their pride together with the spoils of their hands.

extendet manus suas sub eo sicut extendit natans ad natandum et humiliabit gloriam eius cum adlisione manuum eius

25:12 And the fortress of the high fort of thy walls shall 25:12 et munimenta sublimium murorum tuorum concident he bring down, lay low, and bring to the ground, even to the dust.

et humiliabuntur et detrahentur in terram usque ad puluerem

26:1 In that day shall this song be sung in the land of Ju- 26:1 in die illa cantabitur canticum istud in terra Iuda urbs

dah; We have a strong city; salvation will God appoint for walls and bulwarks.

fortitudinis nostrae saluator ponetur in ea murus et antemu-

26:2 Open ye the gates, that the righteous nation which rale 26:2 aperite portas et ingrediatur gens iusta custodiens

keepeth the truth may enter in.

26:3 Thou wilt keep him in perfect peace, whose mind is ueritatem 26:3 uetus error abiit seruabis pacem pacem quia

stayed on thee: because he trusteth in thee.

26:4 Trust ye in the LORD for ever: for in the LORD in te sperauimus 26:4 sperastis in Domino in saeculis aeter-

JEHOVAH is everlasting strength:

Prophetia Isaiae

nis in Domino Deo forti in perpetuum 26:5 quia incuruabit 26:5 For he bringeth down them that dwell on high; the

lofty city, he layeth it low; he layeth it low, even to the ground; he bringeth it even to the dust.

habitantes in excelso ciuitatem sublimem humiliabit humiliabit eam usque ad terram detrahet eam usque ad puluerem 26:6 conculcabit eam pes pedes pauperis gressus egenorum 26:6 The foot shall tread it down, even the feet of the poor,

and the steps of the needy.

26:7 semita iusti recta est rectus callis iusti ad ambulandum 26:7 The way of the just is uprightness: thou, most up-

right, dost weigh the path of the just.

26:8 et in semita iudiciorum tuorum Domine sustinuimus te 26:8 Yea, in the way of thy judgments, O LORD, have we

waited for thee; the desire of our soul is to thy name, and to the remembrance of thee.

nomen tuum et memoriale tuum in desiderio animae 26:9

26:9 With my soul have I desired thee in the night; yea, with my spirit within me will I seek thee early: for when thy judgments are in the earth, the inhabitants of the world will learn righteousness.

anima mea desiderauit te in nocte sed et spiritu meo in praecordiis meis de mane uigilabo ad te cum feceris iudicia tua in terra iustitiam discent habitatores orbis 26:10 misereamur 26:10 Let favour be shewed to the wicked, yet will he not

learn righteousness: in the land of uprightness will he deal unjustly, and will not behold the majesty of the LORD.

impio et non discet iustitiam in terra sanctorum inique gessit et non uidebit gloriam Domini 26:11 Domine exaltetur 26:11 LORD, when thy hand is lifted up, they will not see:

but they shall see, and be ashamed for their envy at the people; yea, the fire of thine enemies shall devour them.

manus tua et non uideant uideant et confundantur zelantes populi et ignis hostes tuos deuoret 26:12 Domine dabis pa- 26:12 LORD, thou wilt ordain peace for us: for thou also

hast wrought all our works in us.

cem nobis omnia enim opera nostra operatus es nobis 26:13 26:13 O LORD our God, other lords beside thee have had

dominion over us: but by thee only will we make mention of thy name.

Domine Deus noster possederunt nos domini absque te tan-

tum in te recordemur nominis tui 26:14 morientes non ui- 26:14 They are dead, they shall not live; they are deceased, they shall not rise: therefore hast thou visited and destroyed them, and made all their memory to perish.

uant gigantes non resurgant propterea uisitasti et contriuisti eos et perdidisti omnem memoriam eorum 26:15 indulsisti 26:15 Thou hast increased the nation, O LORD, thou hast

increased the nation: thou art glorified: thou hadst removed it far unto all the ends of the earth.

genti Domine indulsisti genti numquid glorificatus es elongasti omnes terminos terrae 26:16 Domine in angustia re- 26:16 LORD, in trouble have they visited thee, they

poured out a prayer when thy chastening was upon them.

quisierunt te in tribulatione murmuris doctrina tua eis 26:17 26:17 Like as a woman with child, that draweth near the

time of her delivery, is in pain, and crieth out in her pangs; so have we been in thy sight, O LORD.

sicut quae concipit cum adpropinquauerit ad partum dolens clamat in doloribus suis sic facti sumus a facie tua Domine 26:18 concepimus et quasi parturiuimus et peperimus spiri- 26:18 We have been with child, we have been in pain, we

have as it were brought forth wind; we have not wrought any deliverance in the earth; neither have the inhabitants

tum salutes non fecimus in terra ideo non ceciderunt habita- of the world fallen. tores terrae 26:19 uiuent mortui tui interfecti mei resurgent 26:19 Thy dead men shall live, together with my dead

body shall they arise. Awake and sing, ye that dwell in dust: for thy dew is as the dew of herbs, and the earth

expergiscimini et laudate qui habitatis in puluere quia ros shall cast out the dead. lucis ros tuus et terram gigantum detrahes in ruinam 26:20 26:20 Come, my people, enter thou into thy chambers, and

shut thy doors about thee: hide thyself as it were for a little moment, until the indignation be overpast.

uade populus meus intra in cubicula tua claude ostia tua super te abscondere modicum ad momentum donec pertranseat indignatio 26:21 ecce enim Dominus egreditur de loco suo 26:21 For, behold, the LORD cometh out of his place to

punish the inhabitants of the earth for their iniquity: the earth also shall disclose her blood, and shall no more cover

ut uisitet iniquitatem habitatoris terrae contra eum et reue- her slain.

Biblia Sacra Vulgata

labit terra sanguinem suum et non operiet ultra interfectos suos

27:1 In that day the LORD with his sore and great and 27:1 in die illo uisitabit Dominus in gladio suo duro et

strong sword shall punish leviathan the piercing serpent, even leviathan that crooked serpent; and he shall slay the

dragon that is in the sea. grandi et forti super Leuiathan serpentem uectem et super

Leuiathan serpentem tortuosum et occidet cetum qui in mari

27:2 In that day sing ye unto her, A vineyard of red wine. est 27:2 in die illa uinea meri cantabit ei 27:3 ego Domi-

27:3 I the LORD do keep it; I will water it every moment: lest any hurt it, I will keep it night and day.

nus qui seruo eam repente propinabo ei ne forte uisitetur

27:4 Fury is not in me: who would set the briers and thorns contra eam nocte et die seruo eam 27:4 indignatio non est against me in battle? I would go through them, I would burn them together.

mihi quis dabit me spinam et ueprem in proelio gradiar super

27:5 Or let him take hold of my strength, that he may make eam succendam eam pariter 27:5 an potius tenebit fortitu-

peace with me; and he shall make peace with me. 27:6 He shall cause them that come of Jacob to take root: dinem meam faciet pacem mihi pacem faciet mihi 27:6 qui Israel shall blossom and bud, and fill the face of the world with fruit.

egrediuntur impetu ad Iacob florebit et germinabit Israhel

27:7 Hath he smitten him, as he smote those that smote et implebunt faciem orbis semine 27:7 numquid iuxta pla-

him? or is he slain according to the slaughter of them that are slain by him?

gam percutientis se percussit eum aut sicut occidit interfec-

27:8 In measure, when it shooteth forth, thou wilt debate tos eius sic occisus est 27:8 in mensura contra mensuram with it: he stayeth his rough wind in the day of the east wind.

cum abiecta fuerit iudicabis eam meditata est in spiritu suo

27:9 By this therefore shall the iniquity of Jacob be duro per diem aestus 27:9 idcirco super hoc dimittetur in-

purged; and this is all the fruit to take away his sin; when he maketh all the stones of the altar as chalkstones that are

beaten in sunder, the groves and images shall not stand up. iquitas domui Iacob et iste omnis fructus ut auferatur pecca-

tum eius cum posuerit omnes lapides altaris sicut lapides ci-

27:10 Yet the defenced city shall be desolate, and the habi- neris adlisos non stabunt luci et delubra 27:10 ciuitas enim

tation forsaken, and left like a wilderness: there shall the calf feed, and there shall he lie down, and consume the

branches thereof. munita desolata erit speciosa relinquetur et dimittetur quasi

desertum ibi pascetur uitulus et ibi accubabit et consumet

27:11 When the boughs thereof are withered, they shall be summitates eius 27:11 in siccitate messis illius conterentur

broken off: the women come, and set them on fire: for it is a people of no understanding: therefore he that made them will not have mercy on them, and he that formed them will shew them no favour.

mulieres uenientes et docentes eam non est enim populus sapiens propterea non miserebitur eius qui fecit eum et qui

27:12 And it shall come to pass in that day, that the LORD formauit eum non parcet ei 27:12 et erit in die illa percutiet

shall beat off from the channel of the river unto the stream of Egypt, and ye shall be gathered one by one, O ye chil-

dren of Israel. Dominus ab alueo Fluminis usque ad torrentem Aegypti et 27:13 And it shall come to pass in that day, that the great uos congregabimini unus et unus filii Israhel 27:13 et erit

trumpet shall be blown, and they shall come which were ready to perish in the land of Assyria, and the outcasts in the land of Egypt, and shall worship the LORD in the holy mount at Jerusalem.

in die illa clangetur in tuba magna et uenient qui perditi fuerant de terra Assyriorum et qui eiecti erant in terra Aegypti et adorabunt Dominum in monte sancto in Hierusalem

28:1 Woe to the crown of pride, to the drunkards of 28:1 uae coronae superbiae ebriis Ephraim et flori deci-

Ephraim, whose glorious beauty is a fading flower, which are on the head of the fat valleys of them that are overcome

with wine! denti gloriae exultationis eius qui erant in uertice uallis pin-

Prophetia Isaiae

guissimae errantes a uino 28:2 ecce ualidus et fortis Do- 28:2 Behold, the Lord hath a mighty and strong one, which

as a tempest of hail and a destroying storm, as a flood of mighty waters overflowing, shall cast down to the earth

mini sicut impetus grandinis turbo confringens sicut impetus with the hand. aquarum multarum inundantium et emissarum super terram spatiosam 28:3 pedibus conculcabitur corona superbiae eb- 28:3 The crown of pride, the drunkards of Ephraim, shall

be trodden under feet:

riorum Ephraim 28:4 et erit flos decidens gloriae exultatio- 28:4 And the glorious beauty, which is on the head of the

fat valley, shall be a fading flower, and as the hasty fruit before the summer; which when he that looketh upon it

nis eius qui est super uerticem uallis pinguium quasi tem- seeth, while it is yet in his hand he eateth it up. poraneum ante maturitatem autumni quod cum aspexerit uidens statim ut manu tenuerit deuorabit illud 28:5 in die illa 28:5 In that day shall the LORD of hosts be for a crown of

glory, and for a diadem of beauty, unto the residue of his people,

erit Dominus exercituum corona gloriae et sertum exultationis residuo populi sui 28:6 et spiritus iudicii sedenti super 28:6 And for a spirit of judgment to him that sitteth in

judgment, and for strength to them that turn the battle to the gate.

iudicium et fortitudo reuertentibus de bello ad portam 28:7

28:7 But they also have erred through wine, and through strong drink are out of the way; the priest and the prophet have erred through strong drink, they are swallowed up of wine, they are out of the way through strong drink; they err in vision, they stumble in judgment.

uerum hii quoque prae uino nescierunt et prae ebrietate errauerunt sacerdos et propheta nescierunt prae ebrietate absorti sunt a uino errauerunt in ebrietate nescierunt uidentem igno-

rauerunt iudicium 28:8 omnes enim mensae repletae sunt 28:8 For all tables are full of vomit and filthiness, so that

there is no place clean.

uomitu sordiumque ita ut non esset ultra locus 28:9 quem 28:9 Whom shall he teach knowledge? and whom shall he

make to understand doctrine? them that are weaned from the milk, and drawn from the breasts.

docebit scientiam et quem intellegere faciet auditum ablactatos a lacte apulsos ab uberibus 28:10 quia manda remanda 28:10 For precept must be upon precept, precept upon pre-

cept; line upon line, line upon line; here a little, and there a little:

manda remanda expecta reexpecta expecta reexpecta modicum ibi modicum ibi 28:11 in loquella enim labii et lingua 28:11 For with stammering lips and another tongue will

he speak to this people.

altera loquetur ad populum istum 28:12 cui dixit haec re- 28:12 To whom he said, This is the rest wherewith ye may

cause the weary to rest; and this is the refreshing: yet they would not hear.

quies reficite lassum et hoc est meum refrigerium et noluerunt audire 28:13 et erit eis uerbum Domini manda remanda 28:13 But the word of the LORD was unto them precept

upon precept, precept upon precept; line upon line, line upon line; here a little, and there a little; that they might go, and fall backward, and be broken, and snared, and taken.

manda remanda expecta reexpecta expecta reexpecta modicum ibi modicum ibi ut uadant et cadant retrorsum et conterantur et inlaqueentur et capiantur 28:14 propter hoc audite 28:14 Wherefore hear the word of the LORD, ye scornful

men, that rule this people which is in Jerusalem.

uerbum Domini uiri inlusores qui dominamini super populum meum qui est in Hierusalem 28:15 dixistis enim per- 28:15 Because ye have said, We have made a covenant

with death, and with hell are we at agreement; when the overflowing scourge shall pass through, it shall not come unto us: for we have made lies our refuge, and under falsehood have we hid ourselves:

cussimus foedus cum morte et cum inferno fecimus pactum flagellum inundans cum transierit non ueniet super nos quia posuimus mendacium spem nostram et mendacio protecti sumus 28:16 idcirco haec dicit Dominus Deus ecce ego mit- 28:16 Therefore thus saith the Lord GOD, Behold, I lay

in Zion for a foundation a stone, a tried stone, a precious corner stone, a sure foundation: he that believeth shall not

tam in fundamentis Sion lapidem lapidem probatum angula- make haste.

Biblia Sacra Vulgata

rem pretiosum in fundamento fundatum qui crediderit non

28:17 Judgment also will I lay to the line, and righteous- festinet 28:17 et ponam iudicium in pondere et iustitiam in

ness to the plummet: and the hail shall sweep away the refuge of lies, and the waters shall overflow the hiding

place. mensura et subuertet grando spem mendacii et protectionem

28:18 And your covenant with death shall be disannulled, aquae inundabunt 28:18 et delebitur foedus uestrum cum and your agreement with hell shall not stand; when the overflowing scourge shall pass through, then ye shall be

trodden down by it. morte et pactum uestrum cum inferno non stabit flagellum 28:19 From the time that it goeth forth it shall take you: inundans cum transierit eritis ei in conculcationem 28:19

for morning by morning shall it pass over, by day and by night: and it shall be a vexation only to understand the

report. quandocumque pertransierit tollet uos quoniam mane dilu-

culo pertransibit in die et in nocte et tantummodo sola uexa-

28:20 For the bed is shorter than that a man can stretch tio intellectum dabit auditui 28:20 coangustatum est enim

himself on it: and the covering narrower than that he can wrap himself in it.

stratum ita ut alter decidat et pallium breue utrumque operire

28:21 For the LORD shall rise up as in mount Perazim, he non potest 28:21 sicut enim in monte Diuisionum stabit Do-

shall be wroth as in the valley of Gibeon, that he may do his work, his strange work; and bring to pass his act, his

strange act. minus sicut in ualle quae est in Gabao irascetur ut faciat opus

suum alienum opus eius ut operetur opus suum peregrinum

28:22 Now therefore be ye not mockers, lest your bands be est opus ab eo 28:22 et nunc nolite inludere ne forte conmade strong: for I have heard from the Lord GOD of hosts a consumption, even determined upon the whole earth.

stringantur uincula uestra consummationem enim et adbreuiationem audiui a Domino Deo exercituum super uniuer-

28:23 Give ye ear, and hear my voice; hearken, and hear sam terram 28:23 auribus percipite et audite uocem meam

my speech.

28:24 Doth the plowman plow all day to sow? doth he adtendite et audite eloquium meum 28:24 numquid tota die

open and break the clods of his ground?

28:25 When he hath made plain the face thereof, doth he arabit arans ut serat proscindet et sariet humum suam 28:25

not cast abroad the fitches, and scatter the cummin, and cast in the principal wheat and the appointed barley and

the rie in their place? nonne cum adaequauerit faciem eius seret gith et cyminum

sparget et ponet triticum per ordinem et hordeum et milium

28:26 For his God doth instruct him to discretion, and doth et uiciam in finibus suis 28:26 et erudiet eum illud in iuditeach him. 28:27 For the fitches are not threshed with a threshing in- cio Deus suus docebit eum illud 28:27 non enim in serris strument, neither is a cart wheel turned about upon the cummin; but the fitches are beaten out with a staff, and the

cummin with a rod. triturabitur gith nec rota plaustri super cyminum circumiet 28:28 Bread corn is bruised; because he will not ever be sed in uirga excutietur gith et cyminum in baculo 28:28 pa-

threshing it, nor break it with the wheel of his cart, nor bruise it with his horsemen.

nis autem comminuetur uerum non in perpetuum triturans triturabit illum neque uexabit eum rota plaustri nec in ungu-

28:29 This also cometh forth from the LORD of hosts, lis suis comminuet eum 28:29 et hoc a Domino Deo exer-

which is wonderful in counsel, and excellent in working.

cituum exiuit ut mirabile faceret consilium et magnificaret iustitiam

29:1 Woe to Ariel, to Ariel, the city where David dwelt! 29:1 uae Arihel Arihel ciuitas quam circumdedit Dauid ad-

add ye year to year; let them kill sacrifices. 29:2 Yet I will distress Ariel, and there shall be heaviness ditus est annus ad annum sollemnitates euolutae sunt 29:2 and sorrow: and it shall be unto me as Ariel.

et circumuallabo Arihel et erit tristis et maerens et erit mihi

Prophetia Isaiae

quasi Arihel 29:3 et circumdabo quasi spheram in circuitu 29:3 And I will camp against thee round about, and will

lay siege against thee with a mount, and I will raise forts against thee.

tuo et iaciam contra te aggerem et munimenta ponam in obsidionem tuam 29:4 humiliaberis de terra loqueris et de humo 29:4 And thou shalt be brought down, and shalt speak out

of the ground, and thy speech shall be low out of the dust, and thy voice shall be, as of one that hath a familiar spirit, out of the ground, and thy speech shall whisper out of the dust.

audietur eloquium tuum et erit quasi pythonis de terra uox

tua et de humo eloquium tuum mussitabit 29:5 et erit sicut 29:5 Moreover the multitude of thy strangers shall be like small dust, and the multitude of the terrible ones shall be as chaff that passeth away: yea, it shall be at an instant

puluis tenuis multitudo uentilantium te et sicut fauilla per- suddenly. transiens multitudo eorum qui contra te praeualuerunt 29:6 29:6 Thou shalt be visited of the LORD of hosts with thun-

der, and with earthquake, and great noise, with storm and tempest, and the flame of devouring fire.

eritque repente confestim a Domino exercituum uisitabitur in tonitru et commotione terrae et uoce magna turbinis et

tempestatis et flammae ignis deuorantis 29:7 et erit sic- 29:7 And the multitude of all the nations that fight against Ariel, even all that fight against her and her munition, and that distress her, shall be as a dream of a night vision.

ut somnium uisionis nocturnae multitudo omnium gentium quae dimicauerunt contra Arihel et omnes qui militauerunt et obsederunt et praeualuerunt aduersus eam 29:8 et sicuti 29:8 It shall even be as when an hungry man dreameth,

and, behold, he eateth; but he awaketh, and his soul is empty: or as when a thirsty man dreameth, and, behold, he drinketh; but he awaketh, and, behold, he is faint, and his soul hath appetite: so shall the multitude of all the nations be, that fight against mount Zion.

somniat esuriens et comedit cum autem fuerit expertus uacua est anima eius et sicut somniat sitiens et bibit et postquam fuerit expergefactus lassus adhuc sitit et anima eius uacua est sic erit multitudo omnium gentium quae dimicauerunt contra montem Sion 29:9 obstupescite et admiramini 29:9 Stay yourselves, and wonder; cry ye out, and cry:

they are drunken, but not with wine; they stagger, but not with strong drink.

fluctuate et uacillate inebriamini et non a uino mouemini et non ebrietate 29:10 quoniam miscuit uobis Dominus spi- 29:10 For the LORD hath poured out upon you the spirit

of deep sleep, and hath closed your eyes: the prophets and your rulers, the seers hath he covered.

ritum soporis claudet oculos uestros prophetas et principes uestros qui uident uisiones operiet 29:11 et erit uobis uisio 29:11 And the vision of all is become unto you as the

words of a book that is sealed, which men deliver to one that is learned, saying, Read this, I pray thee: and he saith,

omnium sicut uerba libri signati quem cum dederint scienti I cannot; for it is sealed: litteras dicent lege istum et respondebit non possum signatus est enim 29:12 et dabitur liber nescienti litteras dicetur- 29:12 And the book is delivered to him that is not learned,

saying, Read this, I pray thee: and he saith, I am not learned.

que ei lege et respondebit nescio litteras 29:13 et dixit Do-

29:13 Wherefore the Lord said, Forasmuch as this people draw near me with their mouth, and with their lips do honour me, but have removed their heart far from me, and their fear toward me is taught by the precept of men:

minus eo quod adpropinquat populus iste ore suo et labiis suis glorificat me cor autem eius longe est a me et timuerunt

me mandato hominum et doctrinis 29:14 ideo ecce ego ad- 29:14 Therefore, behold, I will proceed to do a marvellous work among this people, even a marvellous work and a wonder: for the wisdom of their wise men shall perish,

dam ut admirationem faciam populo huic miraculo grandi et and the understanding of their prudent men shall be hid. stupendo peribit enim sapientia a sapientibus eius et intellectus prudentium eius abscondetur 29:15 uae qui profundi 29:15 Woe unto them that seek deep to hide their counsel

from the LORD, and their works are in the dark, and they say, Who seeth us? and who knoweth us?

estis corde ut a Domino abscondatis consilium quorum sunt

Biblia Sacra Vulgata

in tenebris opera et dicunt quis uidet nos et quis nouit nos

29:16 Surely your turning of things upside down shall be 29:16 peruersa est haec uestra cogitatio quasi lutum contra

esteemed as the potter’s clay: for shall the work say of him that made it, He made me not? or shall the thing framed

say of him that framed it, He had no understanding? figulum cogitet et dicat opus factori suo non fecisti me et 29:17 Is it not yet a very little while, and Lebanon shall figmentum dicat fictori suo non intellegis 29:17 nonne ad-

be turned into a fruitful field, and the fruitful field shall be esteemed as a forest?

huc in modico et in breui conuertetur Libanus in Chermel

29:18 And in that day shall the deaf hear the words of the et Chermel in saltum reputabitur 29:18 et audient in die illa

book, and the eyes of the blind shall see out of obscurity, and out of darkness.

surdi uerba libri et de tenebris et caligine oculi caecorum

29:19 The meek also shall increase their joy in the LORD, uidebunt 29:19 et addent mites in Domino laetitiam et pau-

and the poor among men shall rejoice in the Holy One of Israel.

peres homines in Sancto Israhel exultabunt 29:20 quoniam

29:20 For the terrible one is brought to nought, and the scorner is consumed, and all that watch for iniquity are cut off:

defecit qui praeualebat consummatus est inlusor et succisi

29:21 That make a man an offender for a word, and lay a sunt omnes qui uigilabant super iniquitatem 29:21 qui pec-

snare for him that reproveth in the gate, and turn aside the just for a thing of nought.

care faciebant homines in uerbo et arguentem in porta subp-

29:22 Therefore thus saith the LORD, who redeemed lantabant et declinauerunt frustra a iusto 29:22 propter hoc

Abraham, concerning the house of Jacob, Jacob shall not now be ashamed, neither shall his face now wax pale.

haec dicit Dominus ad domum Iacob qui redeeit Abraham non modo confundetur Iacob nec modo uultus eius erube-

29:23 But when he seeth his children, the work of mine scet 29:23 sed cum uiderit filios suos opera manuum me-

hands, in the midst of him, they shall sanctify my name, and sanctify the Holy One of Jacob, and shall fear the God

of Israel. arum in medio sui sanctificantes nomen meum et sanctifi-29:24 They also that erred in spirit shall come to under- cabunt Sanctum Iacob et Deum Israhel praedicabunt 29:24

standing, and they that murmured shall learn doctrine.

et scient errantes spiritu intellectum et mussitatores discent legem

30:1 Woe to the rebellious children, saith the LORD, that 30:1 uae filii desertores dicit Dominus ut faceretis consi-

take counsel, but not of me; and that cover with a covering, but not of my spirit, that they may add sin to sin:

lium et non ex me et ordiremini telam et non per spiritum

30:2 That walk to go down into Egypt, and have not asked meum ut adderetur peccatum super peccatum 30:2 qui am-

at my mouth; to strengthen themselves in the strength of Pharaoh, and to trust in the shadow of Egypt!

bulatis ut descendatis in Aegyptum et os meum non interrogastis sperantes auxilium in fortitudine Pharao et habentes

30:3 Therefore shall the strength of Pharaoh be your fiduciam in umbra Aegypti 30:3 et erit uobis fortitudo Pha-

shame, and the trust in the shadow of Egypt your confusion.

raonis in confusionem et fiducia umbrae Aegypti in igno-

30:4 For his princes were at Zoan, and his ambassadors miniam 30:4 erant enim in Tanis principes tui et nuntii tui

came to Hanes.

30:5 They were all ashamed of a people that could not usque ad Anes peruenerunt 30:5 omnes confusi sunt super

profit them, nor be an help nor profit, but a shame, and also a reproach.

populo qui eis prodeese non potuit non fuerunt in auxilium et in aliquam utilitatem sed in confusionem et obprobrium

30:6 The burden of the beasts of the south: into the land 30:6 onus iumentorum austri in terra tribulationis et angus-

of trouble and anguish, from whence come the young and old lion, the viper and fiery flying serpent, they will carry their riches upon the shoulders of young asses, and their treasures upon the bunches of camels, to a people that shall not profit them.

tiae leaena et leo ex eis uipera et regulus uolans portantes

Prophetia Isaiae

super umeros iumentorum diuitias suas et super gibbum camelorum thesauros suos ad populum qui eis prodeese non poterit 30:7 Aegyptus enim frustra et uane auxiliabitur ideo 30:7 For the Egyptians shall help in vain, and to no pur-

pose: therefore have I cried concerning this, Their strength is to sit still.

clamaui super hoc superbia tantum est quiesce 30:8 nunc

30:8 Now go, write it before them in a table, and note it in a book, that it may be for the time to come for ever and ever:

ingressus scribe eis super buxum et in libro diligenter exara illud et erit in die nouissimo in testimonium usque ad ae-

ternum 30:9 populus enim ad iracundiam prouocans est et 30:9 That this is a rebellious people, lying children, chil-

dren that will not hear the law of the LORD:

filii mendaces filii nolentes audire legem Domini 30:10 qui 30:10 Which say to the seers, See not; and to the prophets,

Prophesy not unto us right things, speak unto us smooth things, prophesy deceits:

dicunt uidentibus nolite uidere et aspicientibus nolite aspicere nobis ea quae recta sunt loquimini nobis placentia uidete nobis errores 30:11 auferte a me uiam declinate a me 30:11 Get you out of the way, turn aside out of the path,

cause the Holy One of Israel to cease from before us.

semitam cesset a facie nostra Sanctus Israhel 30:12 prop- 30:12 Wherefore thus saith the Holy One of Israel, Be-

cause ye despise this word, and trust in oppression and perverseness, and stay thereon:

terea haec dicit Sanctus Israhel pro eo quod reprobastis uerbum hoc et sperastis in calumniam et tumultum et innixi estis super eo 30:13 propterea erit uobis iniquitas haec sicut in- 30:13 Therefore this iniquity shall be to you as a breach

ready to fall, swelling out in a high wall, whose breaking cometh suddenly at an instant.

terruptio cadens et requisita in muro excelso quoniam subito dum non speratur ueniet contritio eius 30:14 et comminue- 30:14 And he shall break it as the breaking of the potters’

vessel that is broken in pieces; he shall not spare: so that there shall not be found in the bursting of it a sherd to take

tur sicut conteritur lagoena figuli contritione perualida et non fire from the hearth, or to take water withal out of the pit. inuenietur de fragmentis eius testa in qua portetur igniculus de incendio aut hauriatur parum aquae de fouea 30:15 quia 30:15 For thus saith the Lord GOD, the Holy One of Israel;

In returning and rest shall ye be saved; in quietness and in confidence shall be your strength: and ye would not.

haec dicit Dominus Deus Sanctus Israhel si reuertamini et quiescatis salui eritis in silentio et in spe erit fortitudo uestra et noluistis 30:16 et dixistis nequaquam sed ad equos fu- 30:16 But ye said, No; for we will flee upon horses; there-

fore shall ye flee: and, We will ride upon the swift; therefore shall they that pursue you be swift.

giemus ideo fugietis et super ueloces ascendemus ideo ueloces erunt qui persequentur uos 30:17 mille homines a fa- 30:17 One thousand shall flee at the rebuke of one; at the

rebuke of five shall ye flee: till ye be left as a beacon upon the top of a mountain, and as an ensign on an hill.

cie terroris unius et a facie terroris quinque fugietis donec relinquamini quasi malus nauis in uertice montis et quasi signum super collem 30:18 propterea expectat Dominus ut 30:18 And therefore will the LORD wait, that he may be

gracious unto you, and therefore will he be exalted, that he may have mercy upon you: for the LORD is a God of

misereatur uestri et ideo exaltabitur parcens uobis quia Deus judgment: blessed are all they that wait for him.

iudicii Dominus beati omnes qui expectant eum 30:19 po- 30:19 For the people shall dwell in Zion at Jerusalem: thou shalt weep no more: he will be very gracious unto thee at the voice of thy cry; when he shall hear it, he will answer

pulus enim Sion habitabit in Hierusalem plorans nequaquam thee. plorabis miserans miserebitur tui ad uocem clamoris tui statim ut audierit respondebit tibi 30:20 et dabit uobis Domi- 30:20 And though the Lord give you the bread of adver-

sity, and the water of affliction, yet shall not thy teachers be removed into a corner any more, but thine eyes shall see thy teachers:

Biblia Sacra Vulgata

nus panem artum et aquam breuem et non faciet auolare a te ultra doctorem tuum et erunt oculi tui uidentes praeceptorem

30:21 And thine ears shall hear a word behind thee, saying, tuum 30:21 et aures tuae audient uerbum post tergum mo-This is the way, walk ye in it, when ye turn to the right hand, and when ye turn to the left.

nentis haec uia ambulate in ea neque ad dexteram neque ad

30:22 Ye shall defile also the covering of thy graven im- sinistram 30:22 et contaminabis lamminas sculptilium ar-

ages of silver, and the ornament of thy molten images of gold: thou shalt cast them away as a menstruous cloth;

thou shalt say unto it, Get thee hence. genti tui et uestimentum conflatilis auri tui et disperges ea 30:23 Then shall he give the rain of thy seed, that thou sicut inmunditiam menstruatae egredere dices ei 30:23 et

shalt sow the ground withal; and bread of the increase of the earth, and it shall be fat and plenteous: in that day shall

thy cattle feed in large pastures. dabitur pluuia semini tuo ubicumque seminaueris in terra et

panis frugum terrae erit uberrimus et pinguis pascetur in pos-

30:24 The oxen likewise and the young asses that ear the sessione tua in die illo agnus spatiose 30:24 et tauri tui et

ground shall eat clean provender, which hath been winnowed with the shovel and with the fan.

pulli asinorum qui operantur terram commixtum migma co-

30:25 And there shall be upon every high mountain, and medent sic in area ut uentilatum est 30:25 et erunt super

upon every high hill, rivers and streams of waters in the day of the great slaughter, when the towers fall.

omnem montem excelsum et super omnem collem eleuatum riui currentium aquarum in die interfectionis multorum cum

30:26 Moreover the light of the moon shall be as the light ceciderint turres 30:26 et erit lux lunae sicut lux solis et lux

of the sun, and the light of the sun shall be sevenfold, as the light of seven days, in the day that the LORD bindeth up the breach of his people, and healeth the stroke of their wound.

solis erit septempliciter sicut lux septem dierum in die qua alligauerit Dominus uulnus populi sui et percussuram pla-

30:27 Behold, the name of the LORD cometh from far, gae eius sanauerit 30:27 ecce nomen Domini uenit de lon-

burning with his anger, and the burden thereof is heavy: his lips are full of indignation, and his tongue as a devour-

ing fire: ginquo ardens furor eius et grauis ad portandum labia eius

repleta sunt indignatione et lingua eius quasi ignis deuorans

30:28 And his breath, as an overflowing stream, shall 30:28 spiritus eius uelut torrens inundans usque ad medium

reach to the midst of the neck, to sift the nations with the sieve of vanity: and there shall be a bridle in the jaws of

the people, causing them to err. colli ad perdendas gentes in nihilum et frenum erroris quod 30:29 Ye shall have a song, as in the night when a holy erat in maxillis populorum 30:29 canticum erit uobis sic-

solemnity is kept; and gladness of heart, as when one goeth with a pipe to come into the mountain of the LORD,

to the mighty One of Israel. ut nox sanctificatae sollemnitatis et laetitia cordis sicut qui

pergit cum tibia ut intret in montem Domini ad Fortem Is-

30:30 And the LORD shall cause his glorious voice to be rahel 30:30 et auditam faciet Dominus gloriam uocis suae

heard, and shall shew the lighting down of his arm, with the indignation of his anger, and with the flame of a devouring fire, with scattering, and tempest, and hailstones. et terrorem brachii sui ostendet in comminatione furoris et

flamma ignis deuorantis adlidet in turbine et in lapide gran-

30:31 For through the voice of the LORD shall the Assyr- dinis 30:31 a uoce enim Domini pauebit Assur uirga perian be beaten down, which smote with a rod.

30:32 And in every place where the grounded staff shall cussus 30:32 et erit transitus uirgae fundatus quam requie-

pass, which the LORD shall lay upon him, it shall be with tabrets and harps: and in battles of shaking will he fight

with it. scere faciet Dominus super eum in tympanis et in citharis 30:33 For Tophet is ordained of old; yea, for the king it is et in bellis praecipuis expugnabit eos 30:33 praeparata est

prepared; he hath made it deep and large: the pile thereof is fire and much wood; the breath of the LORD, like a

stream of brimstone, doth kindle it. enim ab heri Thofeth a rege praeparata profunda et dilatata

Prophetia Isaiae

nutrimenta eius ignis et ligna multa flatus Domini sicut torrens sulphuris succendens eam

31:1 uae qui descendunt in Aegyptum ad auxilium in eq- 31:1 Woe to them that go down to Egypt for help; and stay

on horses, and trust in chariots, because they are many; and in horsemen, because they are very strong; but they look not unto the Holy One of Israel, neither seek the LORD!

uis sperantes et habentes fiduciam super quadrigis quia multae sunt et super equitibus quia praeualidi nimis et non sunt confisi super Sanctum Israhel et Dominum non requisierunt 31:2 ipse autem sapiens adduxit malum et uerba sua non abs- 31:2 Yet he also is wise, and will bring evil, and will not

call back his words: but will arise against the house of the evildoers, and against the help of them that work iniquity.

tulit et consurget contra domum pessimorum et contra auxilium operantium iniquitatem 31:3 Aegyptus homo et non 31:3 Now the Egyptians are men, and not God; and their

horses flesh, and not spirit. When the LORD shall stretch out his hand, both he that helpeth shall fall, and he that is

deus et equi eorum caro et non spiritus et Dominus inclina- holpen shall fall down, and they all shall fail together. bit manum suam et corruet auxiliator et cadet cui praestatur auxilium simulque omnes consumentur 31:4 quia haec di- 31:4 For thus hath the LORD spoken unto me, Like as

the lion and the young lion roaring on his prey, when a multitude of shepherds is called forth against him, he will not be afraid of their voice, nor abase himself for the noise of them: so shall the LORD of hosts come down to fight for mount Zion, and for the hill thereof.

cit Dominus ad me quomodo si rugiat leo et catulus leonis super praedam suam cum occurrerit ei multitudo pastorum a uoce eorum non formidabit et a multitudine eorum non pauebit sic descendet Dominus exercituum ut proelietur super montem Sion et super collem eius 31:5 sicut aues uolantes 31:5 As birds flying, so will the LORD of hosts defend

Jerusalem; defending also he will deliver it; and passing over he will preserve it.

sic proteget Dominus exercituum Hierusalem protegens et liberans transiens et saluans 31:6 conuertimini sicut in pro- 31:6 Turn ye unto him from whom the children of Israel

have deeply revolted.

fundum recesseratis filii Israhel 31:7 in die enim illa abiciet 31:7 For in that day every man shall cast away his idols of

silver, and his idols of gold, which your own hands have made unto you for a sin.

uir idola argenti sui et idola auri sui quae fecerunt uobis manus uestrae in peccatum 31:8 et cadet Assur in gladio non 31:8 Then shall the Assyrian fall with the sword, not of a

mighty man; and the sword, not of a mean man, shall devour him: but he shall flee from the sword, and his young

uiri et gladius non hominis uorabit eum et fugiet non a facie men shall be discomfited. gladii et iuuenes eius uectigales erunt 31:9 et fortitudo eius 31:9 And he shall pass over to his strong hold for fear, and

his princes shall be afraid of the ensign, saith the LORD, whose fire is in Zion, and his furnace in Jerusalem.

a terrore transibit et pauebunt fugientes principes eius dixit Dominus cuius ignis est in Sion et caminus eius in Hierusalem

32:1 ecce in iustitia regnabit rex et principes in iudicio 32:1 Behold, a king shall reign in righteousness, and

princes shall rule in judgment.

praeerunt 32:2 et erit uir sicut qui absconditur a uento et 32:2 And a man shall be as an hiding place from the wind,

and a covert from the tempest; as rivers of water in a dry place, as the shadow of a great rock in a weary land.

celat se a tempestate sicut riui aquarum in siti et umbra petrae prominentis in terra deserta 32:3 non caligabunt oculi 32:3 And the eyes of them that see shall not be dim, and

the ears of them that hear shall hearken.

uidentium et aures audientium diligenter auscultabunt 32:4 32:4 The heart also of the rash shall understand knowl-

edge, and the tongue of the stammerers shall be ready to speak plainly.

et cor stultorum intelleget scientiam et lingua balborum ue-

Biblia Sacra Vulgata

32:5 The vile person shall be no more called liberal, nor lociter loquetur et plane 32:5 non uocabitur ultra is qui in-

the churl said to be bountiful.

sipiens est princeps neque fraudulentus appellabitur maior

32:6 For the vile person will speak villany, and his heart 32:6 stultus enim fatua loquetur et cor eius faciet iniquita-

will work iniquity, to practise hypocrisy, and to utter error against the LORD, to make empty the soul of the hungry,

and he will cause the drink of the thirsty to fail. tem ut perficiat simulationem et loquatur ad Dominum frau-

dulenter et uacuefaciat animam esurientis et potum sitienti

32:7 The instruments also of the churl are evil: he deviseth auferat 32:7 fraudulenti uasa pessima sunt ipse enim cogiwicked devices to destroy the poor with lying words, even when the needy speaketh right.

tationes concinnauit ad perdendos mites in sermone menda-

32:8 But the liberal deviseth liberal things; and by liberal cii cum loqueretur pauper iudicium 32:8 princeps uero ea things shall he stand.

quae digna sunt principe cogitauit et ipse super duces sta-

32:9 Rise up, ye women that are at ease; hear my voice, bit 32:9 mulieres opulentae surgite et audite uocem meam

ye careless daughters; give ear unto my speech.

32:10 Many days and years shall ye be troubled, ye care- filiae confidentes percipite auribus eloquium meum 32:10

less women: for the vintage shall fail, the gathering shall not come.

post dies et annum et uos conturbabimini confidentes con-

32:11 Tremble, ye women that are at ease; be troubled, summata est enim uindemia collectio ultra non ueniet 32:11

ye careless ones: strip you, and make you bare, and gird sackcloth upon your loins.

obstupescite opulentae conturbamini confidentes exuite uos

32:12 They shall lament for the teats, for the pleasant et confundimini accingite lumbos uestros 32:12 super ubera

fields, for the fruitful vine.

32:13 Upon the land of my people shall come up thorns plangite super regione desiderabili super uinea fertili 32:13

and briers; yea, upon all the houses of joy in the joyous city:

super humum populi mei spina et uepres ascendent quanto

32:14 Because the palaces shall be forsaken; the multitude magis super omnes domos gaudii ciuitatis exultantis 32:14 of the city shall be left; the forts and towers shall be for dens for ever, a joy of wild asses, a pasture of flocks;

domus enim dimissa est multitudo urbis relicta est tenebrae et palpatio factae sunt super speluncas usque in aeternum

32:15 Until the spirit be poured upon us from on high, and gaudium onagrorum pascua gregum 32:15 donec effundathe wilderness be a fruitful field, and the fruitful field be counted for a forest.

tur super nos spiritus de excelso et erit desertum in Cher-

32:16 Then judgment shall dwell in the wilderness, and mel et Chermel in saltum reputabitur 32:16 et habitabit in

righteousness remain in the fruitful field.

32:17 And the work of righteousness shall be peace; and solitudine iudicium et iustitia in Chermel sedebit 32:17 et

the effect of righteousness quietness and assurance for ever.

erit opus iustitiae pax et cultus iustitiae silentium et secu-

32:18 And my people shall dwell in a peaceable habita- ritas usque in sempiternum 32:18 et sedebit populus meus

tion, and in sure dwellings, and in quiet resting places;

in pulchritudine pacis et in tabernaculis fiduciae et in requie

32:19 When it shall hail, coming down on the forest; and opulenta 32:19 grando autem in descensione saltus et hu-

the city shall be low in a low place.

32:20 Blessed are ye that sow beside all waters, that send militate humiliabitur ciuitas 32:20 beati qui seminatis super

forth thither the feet of the ox and the ass.

omnes aquas inmittentes pedem bouis et asini

33:1 Woe to thee that spoilest, and thou wast not spoiled; 33:1 uae qui praedaris nonne et ipse praedaberis et qui and dealest treacherously, and they dealt not treacherously with thee! when thou shalt cease to spoil, thou shalt be spoiled; and when thou shalt make an end to deal treacherously, they shall deal treacherously with thee.

spernis nonne et ipse sperneris cum consummaueris depraedationem depraedaberis cum fatigatus desiueris contemnere

Prophetia Isaiae

contemneris 33:2 Domine miserere nostri te expectauimus 33:2 O LORD, be gracious unto us; we have waited for

thee: be thou their arm every morning, our salvation also in the time of trouble.

esto brachium eorum in mane et salus nostra in tempore tribulationis 33:3 a uoce angeli fugerunt populi ab exalta- 33:3 At the noise of the tumult the people fled; at the lift-

ing up of thyself the nations were scattered.

tione tua dispersae sunt gentes 33:4 et congregabuntur spo- 33:4 And your spoil shall be gathered like the gathering of

the caterpiller: as the running to and fro of locusts shall he run upon them.

lia uestra sicut colligitur brucus uelut cum fossae plenae fuerint de eo 33:5 magnificatus est Dominus quoniam habita- 33:5 The LORD is exalted; for he dwelleth on high: he

hath filled Zion with judgment and righteousness.

uit in excelso impleuit Sion iudicio et iustitia 33:6 et erit 33:6 And wisdom and knowledge shall be the stability of

thy times, and strength of salvation: the fear of the LORD is his treasure.

fides in temporibus tuis diuitiae salutis sapientia et scientia timor Domini ipse thesaurus eius 33:7 ecce uidentes clama- 33:7 Behold, their valiant ones shall cry without: the am-

bassadors of peace shall weep bitterly.

bunt foris angeli pacis amare flebunt 33:8 dissipatae sunt 33:8 The highways lie waste, the wayfaring man ceaseth:

he hath broken the covenant, he hath despised the cities, he regardeth no man.

uiae cessauit transiens per semitam irritum factum est pactum proiecit ciuitates non reputauit homines 33:9 luxit et 33:9 The earth mourneth and languisheth: Lebanon is

ashamed and hewn down: Sharon is like a wilderness; and Bashan and Carmel shake off their fruits.

elanguit terra confusus est Libanus et obsorduit et factus est Saron sicut desertum et concussa est Basan et Carmelus 33:10 nunc consurgam dicit Dominus nunc exaltabor 33:10 Now will I rise, saith the LORD; now will I be ex-

alted; now will I lift up myself.

nunc subleuabor 33:11 concipietis ardorem parietis stipu- 33:11 Ye shall conceive chaff, ye shall bring forth stubble:

your breath, as fire, shall devour you.

lam spiritus uester ut ignis uorabit uos 33:12 et erunt populi 33:12 And the people shall be as the burnings of lime: as

thorns cut up shall they be burned in the fire.

quasi de incendio cinis spinae congregatae igni conburentur 33:13 audite qui longe estis quae fecerim et cognoscite uicini 33:13 Hear, ye that are far off, what I have done; and, ye

that are near, acknowledge my might.

fortitudinem meam 33:14 conterriti sunt in Sion peccatores 33:14 The sinners in Zion are afraid; fearfulness hath sur-

prised the hypocrites. Who among us shall dwell with the devouring fire? who among us shall dwell with everlasting

possedit tremor hypocritas quis poterit habitare de uobis cum burnings? igne deuorante quis habitabit ex uobis cum ardoribus sempiternis 33:15 qui ambulat in iustitiis et loquitur ueritates qui 33:15 He that walketh righteously, and speaketh uprightly;

he that despiseth the gain of oppressions, that shaketh his hands from holding of bribes, that stoppeth his ears from

proicit auaritiam ex calumnia et excutit manus suas ab omni hearing of blood, and shutteth his eyes from seeing evil; munere qui obturat aures suas ne audiat sanguinem et claudit oculos suos ne uideat malum 33:16 iste in excelsis habitabit 33:16 He shall dwell on high: his place of defence shall

be the munitions of rocks: bread shall be given him; his waters shall be sure.

munimenta saxorum sublimitas eius panis ei datus est aquae eius fideles sunt 33:17 regem in decore suo uidebunt oculi 33:17 Thine eyes shall see the king in his beauty: they

shall behold the land that is very far off.

eius cernent terram de longe 33:18 cor tuum meditabitur ti- 33:18 Thine heart shall meditate terror. Where is the

scribe? where is the receiver? where is he that counted the towers?

morem ubi est litteratus ubi legis uerba ponderans ubi doctor paruulorum 33:19 populum inpudentem non uidebis popu- 33:19 Thou shalt not see a fierce people, a people of a

deeper speech than thou canst perceive; of a stammering tongue, that thou canst not understand.

lum alti sermonis ita ut non possis intellegere disertitudinem linguae eius in quo nulla est sapientia 33:20 respice Sion 33:20 Look upon Zion, the city of our solemnities: thine

eyes shall see Jerusalem a quiet habitation, a tabernacle that shall not be taken down; not one of the stakes thereof shall ever be removed, neither shall any of the cords thereof be broken.

Biblia Sacra Vulgata

ciuitatem sollemnitatis nostrae oculi tui uidebunt Hierusalem habitationem opulentam tabernaculum quod nequaquam transferri poterit nec auferentur claui eius in sempiternum et

33:21 But there the glorious LORD will be unto us a place omnes funiculi eius non rumpentur 33:21 quia solummodo

of broad rivers and streams; wherein shall go no galley with oars, neither shall gallant ship pass thereby.

ibi magnificus Dominus noster locus fluuiorum riui latissimi et patentes non transibit per eum nauis remigum neque trieris

33:22 For the LORD is our judge, the LORD is our law- magna transgredietur eum 33:22 Dominus enim iudex nos-

giver, the LORD is our king; he will save us.

ter Dominus legifer noster Dominus rex noster ipse saluabit

33:23 Thy tacklings are loosed; they could not well nos 33:23 laxati sunt funiculi tui sed non praeualebunt sic

strengthen their mast, they could not spread the sail: then is the prey of a great spoil divided; the lame take the prey.

erit malus tuus ut dilatare signum non queas tunc diuidentur

33:24 And the inhabitant shall not say, I am sick: the peo- spolia praedarum multarum claudi diripient rapinam 33:24

ple that dwell therein shall be forgiven their iniquity.

nec dicet uicinus elangui populus qui habitat in ea auferetur ab eo iniquitas

34:1 Come near, ye nations, to hear; and hearken, ye peo- 34:1 accedite gentes et audite et populi adtendite audiat

ple: let the earth hear, and all that is therein; the world, and all things that come forth of it.

terra et plenitudo eius orbis et omne germen eius 34:2 quia

34:2 For the indignation of the LORD is upon all nations, and his fury upon all their armies: he hath utterly destroyed them, he hath delivered them to the slaughter.

indignatio Domini super omnes gentes et furor super uniuersam militiam eorum interfecit eos et dedit eos in occisionem

34:3 Their slain also shall be cast out, and their stink shall 34:3 interfecti eorum proicientur et de cadaueribus eorum as-

come up out of their carcases, and the mountains shall be melted with their blood.

cendet fetor tabescent montes sanguine eorum 34:4 et ta-

34:4 And all the host of heaven shall be dissolved, and the heavens shall be rolled together as a scroll: and all their host shall fall down, as the leaf falleth off from the vine, and as a falling fig from the fig tree.

bescet omnis militia caelorum et conplicabuntur sicut liber caeli et omnis militia eorum defluet sicut defluit folium de

34:5 For my sword shall be bathed in heaven: behold, it uinea et de ficu 34:5 quoniam inebriatus est in caelo gla-

shall come down upon Idumea, and upon the people of my curse, to judgment.

dius meus ecce super Idumeam descendet et super populum

34:6 The sword of the LORD is filled with blood, it is interfectionis meae ad iudicium 34:6 gladius Domini rep-

made fat with fatness, and with the blood of lambs and goats, with the fat of the kidneys of rams: for the LORD hath a sacrifice in Bozrah, and a great slaughter in the land of Idumea.

letus est sanguine incrassatus est adipe de sanguine agnorum et hircorum de sanguine medullatorum arietum uictima enim Domini in Bosra et interfectio magna in terra Edom

34:7 And the unicorns shall come down with them, and 34:7 et descendent unicornes cum eis et tauri cum potentibus

the bullocks with the bulls; and their land shall be soaked with blood, and their dust made fat with fatness.

inebriabitur terra eorum sanguine et humus eorum adipe pin-

34:8 For it is the day of the LORD’s vengeance, and the guium 34:8 quia dies ultionis Domini annus retributionum

year of recompences for the controversy of Zion.

34:9 And the streams thereof shall be turned into pitch, iudicii Sion 34:9 et conuertentur torrentes eius in picem et

and the dust thereof into brimstone, and the land thereof shall become burning pitch.

humus eius in sulphur et erit terra eius in picem ardentem

34:10 It shall not be quenched night nor day; the smoke 34:10 nocte et die non extinguetur in sempiternum ascendet

thereof shall go up for ever: from generation to generation it shall lie waste; none shall pass through it for ever and ever.

Prophetia Isaiae

fumus eius a generatione in generationem desolabitur in saeculum saeculorum non erit transiens per eam 34:11 et pos- 34:11 But the cormorant and the bittern shall possess it;

the owl also and the raven shall dwell in it: and he shall stretch out upon it the line of confusion, and the stones of

sidebunt illam onocrotalus et ericius et ibis et coruus habita- emptiness. bunt in ea et extendetur super eam mensura ut rediiatur ad nihilum et perpendiculum in desolationem 34:12 nobiles eius 34:12 They shall call the nobles thereof to the kingdom,

but none shall be there, and all her princes shall be nothing.

non erunt ibi regem potius inuocabunt et omnes principes eius erunt in nihilum 34:13 et orientur in domibus eius spi- 34:13 And thorns shall come up in her palaces, nettles and

brambles in the fortresses thereof: and it shall be an habitation of dragons, and a court for owls.

nae et urticae et paliurus in munitionibus eius et erit cubile draconum et pascua strutionum 34:14 et occurrent daemo- 34:14 The wild beasts of the desert shall also meet with

the wild beasts of the island, and the satyr shall cry to his fellow; the screech owl also shall rest there, and find for

nia onocentauris et pilosus clamabit alter ad alterum ibi cu- herself a place of rest. bauit lamia et inuenit sibi requiem 34:15 ibi habuit foueam 34:15 There shall the great owl make her nest, and lay,

and hatch, and gather under her shadow: there shall the vultures also be gathered, every one with her mate.

ericius et enutriuit catulos et circumfodit et fouit in umbra eius illuc congregati sunt milui alter ad alterum 34:16 re- 34:16 Seek ye out of the book of the LORD, and read: no

one of these shall fail, none shall want her mate: for my mouth it hath commanded, and his spirit it hath gathered

quirite diligenter in libro Domini et legite unum ex eis non them. defuit alter ad alterum non quaesiuit quia quod ex ore meo procedit ille mandauit et spiritus eius ipse congregauit ea 34:17 et ipse misit eis sortem et manus eius diuisit eam illis 34:17 And he hath cast the lot for them, and his hand hath

divided it unto them by line: they shall possess it for ever, from generation to generation shall they dwell therein.

in mensuram usque in aeternum possidebunt eam in generatione et generatione habitabunt in ea

35:1 laetabitur deserta et inuia et exultabit solitudo et flore- 35:1 The wilderness and the solitary place shall be glad

for them; and the desert shall rejoice, and blossom as the rose.

bit quasi lilium 35:2 germinans germinabit et exultabit lae-

35:2 It shall blossom abundantly, and rejoice even with joy and singing: the glory of Lebanon shall be given unto it, the excellency of Carmel and Sharon, they shall see the glory of the LORD, and the excellency of our God.

tabunda et laudans gloria Libani data est ei decor Carmeli et Saron ipsi uidebunt gloriam Domini et decorem Dei nostri 35:3 confortate manus dissolutas et genua debilia roborate 35:3 Strengthen ye the weak hands, and confirm the feeble

knees.

35:4 dicite pusillanimis confortamini nolite timere ecce De- 35:4 Say to them that are of a fearful heart, Be strong, fear

not: behold, your God will come with vengeance, even God with a recompence; he will come and save you.

us uester ultionem adducet retributionis Deus ipse ueniet et saluabit uos 35:5 tunc aperientur oculi caecorum et aures 35:5 Then the eyes of the blind shall be opened, and the

ears of the deaf shall be unstopped.

surdorum patebunt 35:6 tunc saliet sicut ceruus claudus et 35:6 Then shall the lame man leap as an hart, and the

tongue of the dumb sing: for in the wilderness shall waters break out, and streams in the desert.

aperta erit lingua mutorum quia scissae sunt in deserto aquae et torrentes in solitudine 35:7 et quae erat arida in stagnum 35:7 And the parched ground shall become a pool, and the

thirsty land springs of water: in the habitation of dragons, where each lay, shall be grass with reeds and rushes.

et sitiens in fontes aquarum in cubilibus in quibus prius dracones habitabant orietur uiror calami et iunci 35:8 et erit ibi 35:8 And an highway shall be there, and a way, and it shall

be called The way of holiness; the unclean shall not pass over it; but it shall be for those: the wayfaring men, though

semita et uia et uia sancta uocabitur non transibit per eam fools, shall not err therein.

Biblia Sacra Vulgata

pollutus et haec erit nobis directa uia ita ut stulti non errent

35:9 No lion shall be there, nor any ravenous beast shall per eam 35:9 non erit ibi leo et mala bestia non ascendet go up thereon, it shall not be found there; but the redeemed shall walk there:

per eam nec inuenietur ibi et ambulabunt qui liberati fuerint

35:10 And the ransomed of the LORD shall return, and 35:10 et redeepti a Domino conuertentur et uenient in Sion

come to Zion with songs and everlasting joy upon their heads: they shall obtain joy and gladness, and sorrow and

sighing shall flee away. cum laude et laetitia sempiterna super caput eorum gaudium et laetitiam obtinebunt et fugiet dolor et gemitus

36:1 Now it came to pass in the fourteenth year of king 36:1 et factum est in quartodecimo anno regis Ezechiae

Hezekiah, that Sennacherib king of Assyria came up against all the defenced cities of Judah, and took them.

ascendit Sennacherib rex Assyriorum super omnes ciuitates

36:2 And the king of Assyria sent Rabshakeh from Iuda munitas et cepit eas 36:2 et misit rex Assyriorum Rab-

Lachish to Jerusalem unto king Hezekiah with a great army. And he stood by the conduit of the upper pool in

the highway of the fuller’s field. sacen de Lachis in Hierusalem ad regem Ezechiam in manu

graui et stetit in aquaeductu piscinae superioris in uia agri

36:3 Then came forth unto him Eliakim, Hilkiah’s son, Fullonis 36:3 et egressus est ad eum Eliachim filius Helciae

which was over the house, and Shebna the scribe, and Joah, Asaph’s son, the recorder.

qui erat super domum et Sobna scriba et Ioae filius Asaph

36:4 And Rabshakeh said unto them, Say ye now to a commentariis 36:4 et dixit ad eos Rabsaces dicite Eze-

Hezekiah, Thus saith the great king, the king of Assyria, What confidence is this wherein thou trustest?

chiae haec dicit rex magnus rex Assyriorum quae est ista fi-

36:5 I say, sayest thou, (but they are but vain words) I ducia qua confidis 36:5 aut quo consilio uel fortitudine re-

have counsel and strength for war: now on whom dost thou trust, that thou rebellest against me?

bellare disponis super quem habes fiduciam quia recessisti

36:6 Lo, thou trustest in the staff of this broken reed, on a me 36:6 ecce confidis super baculum harundineum con-Egypt; whereon if a man lean, it will go into his hand, and pierce it: so is Pharaoh king of Egypt to all that trust in

him. fractum istum super Aegyptum cui si innisus fuerit homo in-

trabit in manu eius et perforabit eam sic Pharao rex Aegypti

36:7 But if thou say to me, We trust in the LORD our God: omnibus qui confidunt in eo 36:7 quod si responderis mihi

is it not he, whose high places and whose altars Hezekiah hath taken away, and said to Judah and to Jerusalem, Ye

shall worship before this altar? in Domino Deo nostro confidimus nonne ipse est cuius abs-

tulit Ezechias excelsa et altaria et dixit Iudae et Hierusalem

36:8 Now therefore give pledges, I pray thee, to my mas- coram altari isto adorabitis 36:8 et nunc trade te domino

ter the king of Assyria, and I will give thee two thousand horses, if thou be able on thy part to set riders upon them.

meo regi Assyriorum et dabo tibi duo milia equorum nec

36:9 How then wilt thou turn away the face of one captain poteris ex te praebere ascensores eorum 36:9 et quomodo

of the least of my master’s servants, and put thy trust on Egypt for chariots and for horsemen?

sustinebis faciem iudicis unius loci ex seruis domini mei minoribus quod si confidis in Aegypto in quadriga et in equiti-

36:10 And am I now come up without the LORD against bus 36:10 et nunc numquid sine Domino ascendi ad terram

this land to destroy it? the LORD said unto me, Go up against this land, and destroy it.

istam ut disperderem eam Dominus dixit ad me ascende su-

36:11 Then said Eliakim and Shebna and Joah unto Rab- per terram istam et disperde eam 36:11 et dixit Eliachim et

shakeh, Speak, I pray thee, unto thy servants in the Syrian language; for we understand it: and speak not to us in the Jews’ language, in the ears of the people that are on the wall.

Sobna et Ioae ad Rabsacen loquere ad seruos tuos syra lingua intellegimus enim ne loquaris ad nos iudaice in auribus

Prophetia Isaiae

populi qui est super murum 36:12 et dixit ad eos Rabsaces 36:12 But Rabshakeh said, Hath my master sent me to thy

master and to thee to speak these words? hath he not sent me to the men that sit upon the wall, that they may eat

numquid ad dominum tuum et ad te misit me dominus meus their own dung, and drink their own piss with you? ut loquerer omnia uerba ista et non potius ad uiros qui sedent in muro ut comedant stercora sua et bibant urinam pedum suorum uobiscum 36:13 et stetit Rabsaces et clamauit uoce 36:13 Then Rabshakeh stood, and cried with a loud voice

in the Jews’ language, and said, Hear ye the words of the great king, the king of Assyria.

magna iudaice et dixit audite uerba regis magni regis Assyriorum 36:14 haec dicit rex non seducat uos Ezechias quia 36:14 Thus saith the king, Let not Hezekiah deceive you:

for he shall not be able to deliver you.

non poterit eruere uos 36:15 et non uobis tribuat fiduciam 36:15 Neither let Hezekiah make you trust in the LORD,

saying, The LORD will surely deliver us: this city shall not be delivered into the hand of the king of Assyria.

Ezechias super Domino dicens eruens liberabit nos Dominus non dabitur ciuitas ista in manu regis Assyriorum 36:16 36:16 Hearken not to Hezekiah: for thus saith the king of

Assyria, Make an agreement with me by a present, and come out to me: and eat ye every one of his vine, and every one of his fig tree, and drink ye every one the waters of his own cistern;

nolite audire Ezechiam haec enim dicit rex Assyriorum facite mecum benedictionem et egredimini ad me et comedite unusquisque uineam suam et unusquisque ficum suam et bibite unusquisque aquam cisternae suae 36:17 donec ueniam 36:17 Until I come and take you away to a land like your

own land, a land of corn and wine, a land of bread and vineyards.

et tollam uos ad terram quae est ut terra uestra terram frumenti et uini terram panum et uinearum 36:18 ne conturbet 36:18 Beware lest Hezekiah persuade you, saying, the

LORD will deliver us. Hath any of the gods of the nations delivered his land out of the hand of the king of Assyria?

uos Ezechias dicens Dominus liberabit nos numquid liberauerunt dii gentium unusquisque terram suam de manu regis Assyriorum 36:19 ubi est deus Emath et Arfad ubi est deus 36:19 Where are the gods of Hamath and Arphad? where

are the gods of Sepharvaim? and have they delivered Samaria out of my hand?

Seffaruaim numquid liberauerunt Samariam de manu mea 36:20 quis est ex omnibus diis terrarum istarum qui eruerit 36:20 Who are they among all the gods of these lands, that

have delivered their land out of my hand, that the LORD should deliver Jerusalem out of my hand?

terram suam de manu mea ut eruat Dominus Hierusalem de manu mea 36:21 et siluerunt et non responderunt ei uerbum 36:21 But they held their peace, and answered him not a

word: for the king’s commandment was, saying, Answer him not.

mandauerat enim rex dicens ne respondeatis ei 36:22 et in-

36:22 Then came Eliakim, the son of Hilkiah, that was over the household, and Shebna the scribe, and Joah, the son of Asaph, the recorder, to Hezekiah with their clothes rent, and told him the words of Rabshakeh.

gressus est Eliachim filius Helciae qui erat super domum et Sobna scriba et Ioae filius Asaph a commentariis ad Ezechiam scissis uestibus et nuntiauerunt ei uerba Rabsacis

37:1 et factum est cum audisset rex Ezechias scidit uesti- 37:1 And it came to pass, when king Hezekiah heard it,

that he rent his clothes, and covered himself with sackcloth, and went into the house of the LORD.

menta sua et obuolutus est sacco et intrauit in domum Domini 37:2 et misit Eliachim qui erat super domum et Sob- 37:2 And he sent Eliakim, who was over the household,

and Shebna the scribe, and the elders of the priests covered with sackcloth, unto Isaiah the prophet the son of Amoz.

nam scribam et seniores de sacerdotibus opertos saccis ad

Isaiam filium Amos prophetam 37:3 et dixerunt ad eum 37:3 And they said unto him, Thus saith Hezekiah, This day is a day of trouble, and of rebuke, and of blasphemy: for the children are come to the birth, and there is not

haec dicit Ezechias dies tribulationis et correptionis et blas- strength to bring forth.

Biblia Sacra Vulgata

phemiae dies haec quia uenerunt filii usque ad partum et uir-

37:4 It may be the LORD thy God will hear the words tus non est parienti 37:4 si quo modo audiat Dominus Deus

of Rabshakeh, whom the king of Assyria his master hath sent to reproach the living God, and will reprove the words which the LORD thy God hath heard: wherefore lift up thy prayer for the remnant that is left.

tuus uerba Rabsaces quem misit rex Assyriorum dominus suus ad blasphemandum Deum uiuentem et obprobrandum sermonibus quos audiuit Dominus Deus tuus leua ergo ora-

37:5 So the servants of king Hezekiah came to Isaiah. tionem pro reliquiis quae reppertae sunt 37:5 et uenerunt 37:6 And Isaiah said unto them, Thus shall ye say unto serui regis Ezechiae ad Isaiam 37:6 et dixit ad eos Isaias

your master, Thus saith the LORD, Be not afraid of the words that thou hast heard, wherewith the servants of the

king of Assyria have blasphemed me. haec dicetis domino uestro haec dicit Dominus ne timeas a

facie uerborum quae audisti quibus blasphemauerunt pueri

37:7 Behold, I will send a blast upon him, and he shall regis Assyriorum me 37:7 ecce ego dabo ei spiritum et au-

hear a rumour, and return to his own land; and I will cause him to fall by the sword in his own land.

diet nuntium et reuertetur ad terram suam et corruere eum fa-

37:8 So Rabshakeh returned, and found the king of As- ciam gladio in terra sua 37:8 reuersus est autem Rabsaces et

syria warring against Libnah: for he had heard that he was departed from Lachish.

inuenit regem Assyriorum proeliantem aduersus Lobna au-

37:9 And he heard say concerning Tirhakah king of dierat enim quia profectus esset de Lachis 37:9 et audiuit

Ethiopia, He is come forth to make war with thee. And when he heard it, he sent messengers to Hezekiah, saying,

de Tharaca rege Aethiopiae dicentes egressus est ut pugnet contra te quod cum audisset misit nuntios ad Ezechiam di-

37:10 Thus shall ye speak to Hezekiah king of Judah, say- cens 37:10 haec dicetis Ezechiae regi Iudae loquentes non

ing, Let not thy God, in whom thou trustest, deceive thee, saying, Jerusalem shall not be given into the hand of the

king of Assyria. te decipiat Deus tuus in quo tu confidis dicens non dabitur 37:11 Behold, thou hast heard what the kings of Assyria Hierusalem in manu regis Assyriorum 37:11 ecce tu audisti

have done to all lands by destroying them utterly; and shalt thou be delivered?

omnia quae fecerunt reges Assyriorum omnibus terris quas

37:12 Have the gods of the nations delivered them which subuerterunt et tu poteris liberari 37:12 numquid eruerunt

my fathers have destroyed, as Gozan, and Haran, and Rezeph, and the children of Eden which were in Telassar?

eos dii gentium quos subuerterunt patres mei Gozan et Aran

37:13 Where is the king of Hamath, and the king of et Reseph et filios Eden qui erant in Thalassar 37:13 ubi

Arphad, and the king of the city of Sepharvaim, Hena, and Ivah?

est rex Emath et rex Arfad et rex urbis Seffaruaim Anahe

37:14 And Hezekiah received the letter from the hand of et Aua 37:14 et tulit Ezechias libros de manu nuntiorum et

the messengers, and read it: and Hezekiah went up unto the house of the LORD, and spread it before the LORD.

legit eos et ascendit in domum Domini et expandit eos Eze-

37:15 And Hezekiah prayed unto the LORD, saying, chias coram Domino 37:15 et orauit Ezechias ad Dominum 37:16 O LORD of hosts, God of Israel, that dwellest be- dicens 37:16 Domine exercituum Deus Israhel qui sedes su-

tween the cherubims, thou art the God, even thou alone, of all the kingdoms of the earth: thou hast made heaven and

earth. per cherubin tu es Deus solus omnium regnorum terrae tu 37:17 Incline thine ear, O LORD, and hear; open thine fecisti caelum et terram 37:17 inclina Domine aurem tuam

eyes, O LORD, and see: and hear all the words of Sennacherib, which hath sent to reproach the living God.

et audi aperi Domine oculos tuos et uide et audi omnia uerba Sennacherib quae misit ad blasphemandum Deum uiuentem

37:18 Of a truth, LORD, the kings of Assyria have laid 37:18 uere enim Domine desertas fecerunt reges Assyriorum

waste all the nations, and their countries,

Prophetia Isaiae

terras et regiones earum 37:19 et dederunt deos earum igni 37:19 And have cast their gods into the fire: for they were

no gods, but the work of men’s hands, wood and stone: therefore they have destroyed them.

non enim erant dii sed opera manuum hominum lignum et lapis et comminuerunt eos 37:20 et nunc Domine Deus no- 37:20 Now therefore, O LORD our God, save us from his

hand, that all the kingdoms of the earth may know that thou art the LORD, even thou only.

ster salua nos de manu eius et cognoscant omnia regna terrae quia tu es Dominus solus 37:21 et misit Isaias filius Amos 37:21 Then Isaiah the son of Amoz sent unto Hezekiah,

saying, Thus saith the LORD God of Israel, Whereas thou hast prayed to me against Sennacherib king of Assyria:

ad Ezechiam dicens haec dicit Dominus Deus Israhel pro quibus rogasti me de Sennacherib rege Assyriorum 37:22 37:22 This is the word which the LORD hath spoken con-

cerning him; The virgin, the daughter of Zion, hath despised thee, and laughed thee to scorn; the daughter of

hoc est uerbum quod locutus est Dominus super eum de- Jerusalem hath shaken her head at thee. spexit te subsannauit te uirgo filia Sion post te caput mouit filia Hierusalem 37:23 cui exprobrasti et quem blasphemasti 37:23 Whom hast thou reproached and blasphemed? and

against whom hast thou exalted thy voice, and lifted up thine eyes on high? even against the Holy One of Israel.

et super quem exaltasti uocem et leuasti altitudinem oculo-

rum tuorum ad Sanctum Israhel 37:24 in manu seruorum 37:24 By thy servants hast thou reproached the Lord, and hast said, By the multitude of my chariots am I come up to the height of the mountains, to the sides of Lebanon; and I will cut down the tall cedars thereof, and the choice fir trees thereof: and I will enter into the height of his border, and the forest of his Carmel.

tuorum exprobrasti Domino et dixisti in multitudine quadrigarum mearum ego ascendi altitudinem montium iuga Libani et succidam excelsa cedrorum eius electas abietes illius et introibo altitudinem summitatis eius saltum Carmeli eius 37:25 ego fodi et bibi aquam et exsiccaui uestigio pedis mei 37:25 I have digged, and drunk water; and with the sole

of my feet have I dried up all the rivers of the besieged places.

omnes riuos aggerum 37:26 numquid non audisti quae olim

37:26 Hast thou not heard long ago, how I have done it; and of ancient times, that I have formed it? now have I brought it to pass, that thou shouldest be to lay waste defenced cities into ruinous heaps.

fecerim ei ex diebus antiquis ego plasmaui illud et nunc adduxi et factum est in eradicationem collium conpugnantium et ciuitatum munitarum 37:27 habitatores earum breuiata 37:27 Therefore their inhabitants were of small power,

they were dismayed and confounded: they were as the grass of the field, and as the green herb, as the grass on

manu contremuerunt et confusi sunt facti sunt sicut faenum the housetops, and as corn blasted before it be grown up. agri et gramen pascuae et herba tectorum quae exaruit antequam maturesceret 37:28 habitationem tuam et egressum 37:28 But I know thy abode, and thy going out, and thy

coming in, and thy rage against me.

tuum et introitum tuum cognoui et insaniam tuam contra me 37:29 cum fureres aduersum me superbia tua ascendit in au- 37:29 Because thy rage against me, and thy tumult, is

come up into mine ears, therefore will I put my hook in thy nose, and my bridle in thy lips, and I will turn thee

res meas ponam ergo circulum in naribus tuis et frenum in back by the way by which thou camest. labiis tuis et reducam te in uiam per quam uenisti 37:30 tibi 37:30 And this shall be a sign unto thee, Ye shall eat this

year such as groweth of itself; and the second year that which springeth of the same: and in the third year sow ye,

autem hoc erit signum comede hoc anno quae sponte nas- and reap, and plant vineyards, and eat the fruit thereof.

cuntur et in anno secundo pomis uescere in anno autem tertio seminate et metite et plantate uineas et comedite fructum earum 37:31 et mittet id quod saluatum fuerit de domo Iuda 37:31 And the remnant that is escaped of the house of Ju-

dah shall again take root downward, and bear fruit upward:

et quod reliquum est radicem deorsum et faciet fructum sur-

Biblia Sacra Vulgata

37:32 For out of Jerusalem shall go forth a remnant, and sum 37:32 quia de Hierusalem exibunt reliquiae et saluatio

they that escape out of mount Zion: the zeal of the LORD of hosts shall do this.

de monte Sion zelus Domini exercituum faciet istud 37:33

37:33 Therefore thus saith the LORD concerning the king of Assyria, He shall not come into this city, nor shoot an arrow there, nor come before it with shields, nor cast a bank against it.

propterea haec dicit Dominus de rege Assyriorum non introibit ciuitatem hanc et non iaciet ibi sagittam et non occupabit

37:34 By the way that he came, by the same shall he re- eam clypeus et non mittet in circuitu eius aggerem 37:34 in

turn, and shall not come into this city, saith the LORD.

uia qua uenit per eam reuertetur et ciuitatem hanc non in-

37:35 For I will defend this city to save it for mine own gredietur dicit Dominus 37:35 et protegam ciuitatem istam

sake, and for my servant David’s sake.

ut saluem eam propter me et propter Dauid seruum meum

37:36 Then the angel of the LORD went forth, and smote 37:36 egressus est autem angelus Domini et percussit in ca-

in the camp of the Assyrians a hundred and fourscore and five thousand: and when they arose early in the morning,

behold, they were all dead corpses. stris Assyriorum centum octoginta quinque milia et surrexe-

37:37 So Sennacherib king of Assyria departed, and went runt mane et ecce omnes cadauera mortuorum 37:37 et egand returned, and dwelt at Nineveh.

ressus est et abiit et reuersus est Sennacherib rex Assyriorum

37:38 And it came to pass, as he was worshipping in the et habitauit in Nineue 37:38 et factum est cum adoraret in

house of Nisroch his god, that Adrammelech and Sharezer his sons smote him with the sword; and they escaped into the land of Armenia: and Esarhaddon his son reigned in his stead.

templo Nesrach deum suum Adramelech et Sarasar filii eius percusserunt eum gladio fugeruntque in terram Ararat et regnauit Asoraddon filius eius pro eo

38:1 In those days was Hezekiah sick unto death. And 38:1 in diebus illis aegrotauit Ezechias usque ad mortem et

Isaiah the prophet the son of Amoz came unto him, and said unto him, Thus saith the LORD, Set thine house in

order: for thou shalt die, and not live. introiuit ad eum Isaias filius Amos propheta et dixit ei haec

dicit Dominus dispone domui tuae quia morieris tu et non

38:2 Then Hezekiah turned his face toward the wall, and uiues 38:2 et conuertit Ezechias faciem suam ad parietem

prayed unto the LORD,

38:3 And said, Remember now, O LORD, I beseech thee, et orauit ad Dominum 38:3 et dixit obsecro Domine me-

how I have walked before thee in truth and with a perfect heart, and have done that which is good in thy sight. And

Hezekiah wept sore. mento quaeso quomodo ambulauerim coram te in ueritate et

in corde perfecto et quod bonum est in oculis tuis fecerim et

38:4 Then came the word of the LORD to Isaiah, saying, fleuit Ezechias fletu magno 38:4 et factum est uerbum Do-38:5 Go, and say to Hezekiah, Thus saith the LORD, the mini ad Isaiam dicens 38:5 uade et dic Ezechiae haec dicit

God of David thy father, I have heard thy prayer, I have seen thy tears: behold, I will add unto thy days fifteen

years. Dominus Deus Dauid patris tui audiui orationem tuam uidi

lacrimam tuam ecce ego adiciam super dies tuos quindecim

38:6 And I will deliver thee and this city out of the hand annos 38:6 et de manu regis Assyriorum eruam te et ciuita-

of the king of Assyria: and I will defend this city.

38:7 And this shall be a sign unto thee from the LORD, tem istam et protegam eam 38:7 hoc autem tibi erit signum

that the LORD will do this thing that he hath spoken;

a Domino quia faciet Dominus uerbum hoc quod locutus est

38:8 Behold, I will bring again the shadow of the degrees, 38:8 ecce ego reuerti faciam umbram linearum per quas dewhich is gone down in the sun dial of Ahaz, ten degrees backward. So the sun returned ten degrees, by which de-

grees it was gone down. scenderat in horologio Ahaz in sole retrorsum decem lineis et reuersus est sol decem lineis per gradus quos descenderat

Prophetia Isaiae

38:9 scriptura Ezechiae regis Iuda cum aegrotasset et con- 38:9 The writing of Hezekiah king of Judah, when he had

been sick, and was recovered of his sickness:

ualuisset de infirmitate sua 38:10 ego dixi in dimidio die- 38:10 I said in the cutting off of my days, I shall go to

the gates of the grave: I am deprived of the residue of my years.

rum meorum uadam ad portas inferi quaesiui residuum an-

norum meorum 38:11 dixi non uidebo Dominum Dominum 38:11 I said, I shall not see the LORD, even the LORD, in the land of the living: I shall behold man no more with the inhabitants of the world.

in terra uiuentium non aspiciam hominem ultra et habitatorem quieuit 38:12 generatio mea ablata est et conuoluta 38:12 Mine age is departed, and is removed from me as a

shepherd’s tent: I have cut off like a weaver my life: he will cut me off with pining sickness: from day even to

est a me quasi tabernaculum pastorum praecisa est uelut a night wilt thou make an end of me. texente uita mea dum adhuc ordirer succidit me de mane us-

que ad uesperam finies me 38:13 sperabam usque ad mane 38:13 I reckoned till morning, that, as a lion, so will he break all my bones: from day even to night wilt thou make an end of me.

quasi leo sic contriuit omnia ossa mea de mane usque ad uesperam finies me 38:14 sicut pullus hirundinis sic clamabo 38:14 Like a crane or a swallow, so did I chatter: I did

mourn as a dove: mine eyes fail with looking upward: O LORD, I am oppressed; undertake for me.

meditabor ut columba adtenuati sunt oculi mei suspicientes in excelsum Domine uim patior sponde pro me 38:15 quid 38:15 What shall I say? he hath both spoken unto me, and

himself hath done it: I shall go softly all my years in the bitterness of my soul.

dicam aut quid respondebit mihi cum ipse fecerit recogitabo omnes annos meos in amaritudine animae meae 38:16 Do- 38:16 O LORD, by these things men live, and in all these

things is the life of my spirit: so wilt thou recover me, and make me to live.

mine sic uiuitur et in talibus uita spiritus mei corripies me et uiuificabis me 38:17 ecce in pace amaritudo mea amaris- 38:17 Behold, for peace I had great bitterness: but thou

hast in love to my soul delivered it from the pit of corruption: for thou hast cast all my sins behind thy back.

sima tu autem eruisti animam meam ut non periret proiecisti

post tergum tuum omnia peccata mea 38:18 quia non infer- 38:18 For the grave cannot praise thee, death can not celebrate thee: they that go down into the pit cannot hope for thy truth.

nus confitebitur tibi neque mors laudabit te non expectabunt

qui descendunt in lacum ueritatem tuam 38:19 uiuens ui- 38:19 The living, the living, he shall praise thee, as I do this day: the father to the children shall make known thy truth.

uens ipse confitebitur tibi sicut et ego hodie pater filiis notam faciet ueritatem tuam 38:20 Domine saluum me fac et 38:20 The LORD was ready to save me: therefore we will

sing my songs to the stringed instruments all the days of our life in the house of the LORD.

psalmos nostros cantabimus cunctis diebus uitae nostrae in domo Domini 38:21 et iussit Isaias ut tollerent massam de 38:21 For Isaiah had said, Let them take a lump of figs,

and lay it for a plaister upon the boil, and he shall recover.

ficis et cataplasmarent super uulnus et sanaretur 38:22 et 38:22 Hezekiah also had said, What is the sign that I shall

go up to the house of the LORD?

dixit Ezechias quod erit signum quia ascendam in domo Domini

39:1 in tempore illo misit Marodach Baladan filius Bala- 39:1 At that time Merodachbaladan, the son of Baladan,

king of Babylon, sent letters and a present to Hezekiah: for he had heard that he had been sick, and was recovered.

dan rex Babylonis libros et munera ad Ezechiam audierat enim quod aegrotasset et conualuisset 39:2 laetatus est au- 39:2 And Hezekiah was glad of them, and shewed them

the house of his precious things, the silver, and the gold, and the spices, and the precious ointment, and all the house of his armour, and all that was found in his treasures: there was nothing in his house, nor in all his dominion, that Hezekiah shewed them not.

tem super eis Ezechias et ostendit eis cellam aromatum et argenti et auri et odoramentorum et unguenti optimi et omnes

Biblia Sacra Vulgata

apothecas supellectilis suae et uniuersa quae inuenta sunt in thesauris eius non fuit uerbum quod non ostenderet eis Eze-

39:3 Then came Isaiah the prophet unto king Hezekiah, chias in domo sua et in omni potestate sua 39:3 introiit au-

and said unto him, What said these men? and from whence came they unto thee? And Hezekiah said, They are come

from a far country unto me, even from Babylon. tem Isaias propheta ad regem Ezechiam et dixit ei quid dixe-

runt uiri isti et unde uenerunt ad te et dixit Ezechias de terra

39:4 Then said he, What have they seen in thine house? longinqua uenerunt ad me de Babylone 39:4 et dixit quid

And Hezekiah answered, All that is in mine house have they seen: there is nothing among my treasures that I have

not shewed them. uiderunt in domo tua et dixit Ezechias omnia quae in domo

mea sunt uiderunt non fuit res quam non ostenderim eis in

39:5 Then said Isaiah to Hezekiah, Hear the word of the thesauris meis 39:5 et dixit Isaias ad Ezechiam audi uer-

LORD of hosts: 39:6 Behold, the days come, that all that is in thine house, bum Domini exercituum 39:6 ecce dies uenient et auferenand that which thy fathers have laid up in store until this day, shall be carried to Babylon: nothing shall be left, saith

the LORD. tur omnia quae in domo tua sunt et quae thesaurizauerunt

patres tui usque ad diem hanc in Babylonem non relinquetur

39:7 And of thy sons that shall issue from thee, which thou quicquam dicit Dominus 39:7 et de filiis tuis qui exibunt shalt beget, shall they take away; and they shall be eunuchs in the palace of the king of Babylon.

de te quos genueris tollent et erunt eunuchi in palatio regis

39:8 Then said Hezekiah to Isaiah, Good is the word of Babylonis 39:8 et dixit Ezechias ad Isaiam bonum uerbum

the LORD which thou hast spoken. He said moreover, For there shall be peace and truth in my days.

Domini quod locutus est et dixit fiat tantum pax et ueritas in diebus meis

40:1 Comfort ye, comfort ye my people, saith your God. 40:1 consolamini consolamini populus meus dicit Deus

40:2 Speak ye comfortably to Jerusalem, and cry unto her, uester 40:2 loquimini ad cor Hierusalem et auocate eam that her warfare is accomplished, that her iniquity is pardoned: for she hath received of the LORD’s hand double

for all her sins. quoniam conpleta est malitia eius dimissa est iniquitas il-

lius suscepit de manu Domini duplicia pro omnibus peccatis

40:3 The voice of him that crieth in the wilderness, Pre- suis 40:3 uox clamantis in deserto parate uiam Domini re-

pare ye the way of the LORD, make straight in the desert a highway for our God.

ctas facite in solitudine semitas Dei nostri 40:4 omnis ual-

40:4 Every valley shall be exalted, and every mountain and hill shall be made low: and the crooked shall be made straight, and the rough places plain:

lis exaltabitur et omnis mons et collis humiliabitur et erunt

40:5 And the glory of the LORD shall be revealed, and all praua in directa et aspera in uias planas 40:5 et reuelabitur

flesh shall see it together: for the mouth of the LORD hath spoken it.

gloria Domini et uidebit omnis caro pariter quod os Domini

40:6 The voice said, Cry. And he said, What shall I cry? locutum est 40:6 uox dicentis clama et dixi quid clamabo

All flesh is grass, and all the goodliness thereof is as the flower of the field:

omnis caro faenum et omnis gloria eius quasi flos agri 40:7

40:7 The grass withereth, the flower fadeth: because the spirit of the LORD bloweth upon it: surely the people is grass.

exsiccatum est faenum et cecidit flos quia spiritus Domini

40:8 The grass withereth, the flower fadeth: but the word sufflauit in eo uere faenum est populus 40:8 exsiccatum est

of our God shall stand for ever.

faenum cecidit flos uerbum autem Dei nostri stabit in aeter-

40:9 O Zion, that bringest good tidings, get thee up into the num 40:9 super montem excelsum ascende tu quae euangehigh mountain; O Jerusalem, that bringest good tidings, lift up thy voice with strength; lift it up, be not afraid; say

unto the cities of Judah, Behold your God! lizas Sion exalta in fortitudine uocem tuam quae euangelizas

Prophetia Isaiae

Hierusalem exalta noli timere dic ciuitatibus Iudae ecce Deus uester 40:10 ecce Dominus Deus in fortitudine ueniet et 40:10 Behold, the Lord GOD will come with strong hand,

and his arm shall rule for him: behold, his reward is with him, and his work before him.

brachium eius dominabitur ecce merces eius cum eo et opus illius coram eo 40:11 sicut pastor gregem suum pascet in 40:11 He shall feed his flock like a shepherd: he shall

gather the lambs with his arm, and carry them in his bosom, and shall gently lead those that are with young.

brachio suo congregabit agnos et in sinu suo leuabit fetas ipse portabit 40:12 quis mensus est pugillo aquas et caelos 40:12 Who hath measured the waters in the hollow of his

hand, and meted out heaven with the span, and comprehended the dust of the earth in a measure, and weighed the

palmo ponderauit quis adpendit tribus digitis molem terrae mountains in scales, and the hills in a balance? et librauit in pondere montes et colles in statera 40:13 quis 40:13 Who hath directed the Spirit of the LORD, or being

his counsellor hath taught him?

adiuuit spiritum Domini aut quis consiliarius eius fuit et ostendit illi 40:14 cum quo iniit consilium et instruxit eum et 40:14 With whom took he counsel, and who instructed

him, and taught him in the path of judgment, and taught him knowledge, and shewed to him the way of understand-

docuit eum semitam iustitiae et erudiuit eum scientiam et ing? uiam prudentiae ostendit illi 40:15 ecce gentes quasi stilla 40:15 Behold, the nations are as a drop of a bucket, and

are counted as the small dust of the balance: behold, he taketh up the isles as a very little thing.

situlae et quasi momentum staterae reputatae sunt ecce insulae quasi puluis exiguus 40:16 et Libanus non sufficiet 40:16 And Lebanon is not sufficient to burn, nor the beasts

thereof sufficient for a burnt offering.

ad succendendum et animalia eius non sufficient ad holo-

caustum 40:17 omnes gentes quasi non sint sic sunt co- 40:17 All nations before him are as nothing; and they are

counted to him less than nothing, and vanity.

ram eo et quasi nihilum et inane reputatae sunt ei 40:18 cui 40:18 To whom then will ye liken God? or what likeness

will ye compare unto him?

ergo similem fecistis Deum aut quam imaginem ponetis ei 40:19 numquid sculptile conflauit faber aut aurifex auro fi- 40:19 The workman melteth a graven image, and the gold-

smith spreadeth it over with gold, and casteth silver chains.

gurauit illud et lamminis argenteis argentarius 40:20 forte 40:20 He that is so impoverished that he hath no oblation

chooseth a tree that will not rot; he seeketh unto him a cunning workman to prepare a graven image, that shall

lignum et inputribile elegit artifex sapiens quaerit quomodo not be moved. statuat simulacrum quod non moueatur 40:21 numquid non 40:21 Have ye not known? have ye not heard? hath it not

been told you from the beginning? have ye not understood from the foundations of the earth?

scietis numquid non audietis numquid non adnuntiatum est ab initio uobis numquid non intellexistis fundamenta terrae 40:22 qui sedet super gyrum terrae et habitatores eius sunt 40:22 It is he that sitteth upon the circle of the earth, and

the inhabitants thereof are as grasshoppers; that stretcheth out the heavens as a curtain, and spreadeth them out as a

quasi lucustae qui extendit uelut nihilum caelos et expandit tent to dwell in: eos sicut tabernaculum ad inhabitandum 40:23 qui dat sec- 40:23 That bringeth the princes to nothing; he maketh the

judges of the earth as vanity.

retorum scrutatores quasi non sint iudices terrae uelut inane fecit 40:24 et quidem neque plantatos neque satos neque ra- 40:24 Yea, they shall not be planted; yea, they shall not be

sown: yea, their stock shall not take root in the earth: and he shall also blow upon them, and they shall wither, and

dicato in terra trunco eorum repente flauit in eos et aruerunt the whirlwind shall take them away as stubble. et turbo quasi stipulam auferet eos 40:25 et cui adsimilastis 40:25 To whom then will ye liken me, or shall I be equal?

saith the Holy One.

me et adaequastis dicit Sanctus 40:26 leuate in excelsum 40:26 Lift up your eyes on high, and behold who hath cre-

ated these things, that bringeth out their host by number: he calleth them all by names by the greatness of his might,

oculos uestros et uidete quis creauit haec qui educit in nu- for that he is strong in power; not one faileth.

Biblia Sacra Vulgata

mero militiam eorum et omnes ex nomine uocat prae multitudine fortitudinis et roboris uirtutisque eius neque unum

40:27 Why sayest thou, O Jacob, and speakest, O Israel, reliquum fuit 40:27 quare dicis Iacob et loqueris Israhel ab-

My way is hid from the LORD, and my judgment is passed over from my God?

scondita est uia mea a Domino et a Deo meo iudicium meum

40:28 Hast thou not known? hast thou not heard, that the transibit 40:28 numquid nescis aut non audisti Deus sempi-

everlasting God, the LORD, the Creator of the ends of the earth, fainteth not, neither is weary? there is no searching

of his understanding. ternus Dominus qui creauit terminos terrae non deficiet ne-40:29 He giveth power to the faint; and to them that have que laborabit nec est inuestigatio sapientiae eius 40:29 qui

no might he increaseth strength.

dat lasso uirtutem et his qui non sunt fortitudinem et robur

40:30 Even the youths shall faint and be weary, and the multiplicat 40:30 deficient pueri et laborabunt et iuuenes in

young men shall utterly fall:

40:31 But they that wait upon the LORD shall renew their infirmitate cadent 40:31 qui autem sperant in Domino mu-

strength; they shall mount up with wings as eagles; they shall run, and not be weary; and they shall walk, and not

faint. tabunt fortitudinem adsument pinnas sicut aquilae current et non laborabunt ambulabunt et non deficient

41:1 Keep silence before me, O islands; and let the people 41:1 taceant ad me insulae et gentes mutent fortitudinem renew their strength: let them come near; then let them speak: let us come near together to judgment.

accedant et tunc loquantur simul ad iudicium propinquemus

41:2 Who raised up the righteous man from the east, called 41:2 quis suscitauit ab oriente iustum uocauit eum ut seque-

him to his foot, gave the nations before him, and made him rule over kings? he gave them as the dust to his sword, and

as driven stubble to his bow. retur se dabit in conspectu eius gentes et reges obtinebit da-

bit quasi puluerem gladio eius sicut stipulam uento raptam

41:3 He pursued them, and passed safely; even by the way arcui eius 41:3 persequetur eos transibit in pace semita in

that he had not gone with his feet.

41:4 Who hath wrought and done it, calling the genera- pedibus eius non apparebit 41:4 quis haec operatus est et

tions from the beginning? I the LORD, the first, and with the last; I am he.

fecit uocans generationes ab exordio ego Dominus primus et

41:5 The isles saw it, and feared; the ends of the earth were nouissimus ego sum 41:5 uiderunt insulae et timuerunt exafraid, drew near, and came.

trema terrae obstipuerunt adpropinquauerunt et accesserunt

41:6 They helped every one his neighbour; and every one 41:6 unusquisque proximo suo auxiliatur et fratri suo dicit

said to his brother, Be of good courage.

41:7 So the carpenter encouraged the goldsmith, and he confortare 41:7 confortabit faber aerarius percutiens mal-

that smootheth with the hammer him that smote the anvil, saying, It is ready for the sodering: and he fastened it with

nails, that it should not be moved. leo eum qui cudebat tunc temporis dicens glutino bonum est 41:8 But thou, Israel, art my servant, Jacob whom I have et confortauit eum in clauis ut non moueatur 41:8 et tu Is-

chosen, the seed of Abraham my friend.

rahel serue meus Iacob quem elegi semen Abraham amici

41:9 Thou whom I have taken from the ends of the earth, mei 41:9 in quo adprehendi te ab extremis terrae et a lon-

and called thee from the chief men thereof, and said unto thee, Thou art my servant; I have chosen thee, and not cast

thee away. ginquis eius uocaui te et dixi tibi seruus meus es tu elegi te et 41:10 Fear thou not; for I am with thee: be not dismayed; non abieci te 41:10 ne timeas quia tecum sum ego ne decli-

for I am thy God: I will strengthen thee; yea, I will help thee; yea, I will uphold thee with the right hand of my

righteousness. nes quia ego Deus tuus confortaui te et auxiliatus sum tui et 41:11 Behold, all they that were incensed against thee suscepi te dextera iusti mei 41:11 ecce confundentur et eru-

shall be ashamed and confounded: they shall be as nothing; and they that strive with thee shall perish.

bescent omnes qui pugnant aduersum te erunt quasi non sint

Prophetia Isaiae

et peribunt uiri qui contradicunt tibi 41:12 quaeres eos et 41:12 Thou shalt seek them, and shalt not find them, even

them that contended with thee: they that war against thee shall be as nothing, and as a thing of nought.

non inuenies uiros rebelles tuos erunt quasi non sint et ueluti consumptio homines bellantes aduersum te 41:13 quia ego 41:13 For I the LORD thy God will hold thy right hand,

saying unto thee, Fear not; I will help thee.

Dominus Deus tuus adprehendens manum tuam dicensque tibi ne timeas ego adiuui te 41:14 noli timere uermis Iacob 41:14 Fear not, thou worm Jacob, and ye men of Israel;

I will help thee, saith the LORD, and thy redeemer, the Holy One of Israel.

qui mortui estis ex Israhel ego auxiliatus sum tui dicit Dominus et redeeptor tuus Sanctus Israhel 41:15 ego posui te 41:15 Behold, I will make thee a new sharp threshing in-

strument having teeth: thou shalt thresh the mountains, and beat them small, and shalt make the hills as chaff.

quasi plaustrum triturans nouum habens rostra serrantia triturabis montes et comminues et colles quasi puluerem pones 41:16 uentilabis eos et uentus tollet et turbo disperget eos et 41:16 Thou shalt fan them, and the wind shall carry them

away, and the whirlwind shall scatter them: and thou shalt rejoice in the LORD, and shalt glory in the Holy One of

tu exultabis in Domino in Sancto Israhel laetaberis 41:17 Israel.

41:17 When the poor and needy seek water, and there is none, and their tongue faileth for thirst, I the LORD will hear them, I the God of Israel will not forsake them.

egeni et pauperes quaerunt aquas et non sunt lingua eorum siti aruit ego Dominus exaudiam eos Deus Israhel non derelinquam eos 41:18 aperiam in supinis collibus flumina et in 41:18 I will open rivers in high places, and fountains in

the midst of the valleys: I will make the wilderness a pool of water, and the dry land springs of water.

medio camporum fontes ponam desertum in stagna aquarum et terram inuiam in riuos aquarum 41:19 dabo in solitudine 41:19 I will plant in the wilderness the cedar, the shittah

tree, and the myrtle, and the oil tree; I will set in the desert the fir tree, and the pine, and the box tree together:

cedrum et spinam et myrtum et lignum oliuae ponam in de-

serto abietem ulmum et buxum simul 41:20 ut uideant et 41:20 That they may see, and know, and consider, and understand together, that the hand of the LORD hath done this, and the Holy One of Israel hath created it.

sciant et recogitent et intellegant pariter quia manus Domini fecit hoc et Sanctus Israhel creauit illud 41:21 prope facite 41:21 Produce your cause, saith the LORD; bring forth

your strong reasons, saith the King of Jacob.

iudicium uestrum dicit Dominus adferte si quid forte habetis dixit Rex Iacob 41:22 accedant et nuntient nobis quaecum- 41:22 Let them bring them forth, and shew us what shall

happen: let them shew the former things, what they be, that we may consider them, and know the latter end of

que uentura sunt priora quae fuerint nuntiate et ponemus cor them; or declare us things for to come. nostrum et sciemus nouissima eorum et quae uentura sunt indicate nobis 41:23 adnuntiate quae uentura sunt in futurum 41:23 Shew the things that are to come hereafter, that we

may know that ye are gods: yea, do good, or do evil, that we may be dismayed, and behold it together.

et sciemus quia dii estis uos bene quoque aut male si potestis facite et loquamur et uideamus simul 41:24 ecce uos estis 41:24 Behold, ye are of nothing, and your work of nought:

an abomination is he that chooseth you.

ex nihilo et opus uestrum ex eo quod non est abominatio est qui elegit uos 41:25 suscitaui ab aquilone et uenit ab ortu 41:25 I have raised up one from the north, and he shall

come: from the rising of the sun shall he call upon my name: and he shall come upon princes as upon morter,

solis uocabit nomen meum et adducet magistratus quasi lu- and as the potter treadeth clay. tum et uelut plastes conculcans humum 41:26 quis adnun- 41:26 Who hath declared from the beginning, that we may

know? and beforetime, that we may say, He is righteous? yea, there is none that sheweth, yea, there is none that de-

tiauit ab exordio ut sciamus et a principio ut dicamus iustus clareth, yea, there is none that heareth your words. es non est neque adnuntians neque praedicens neque audiens

Biblia Sacra Vulgata

41:27 The first shall say to Zion, Behold, behold them: and sermones uestros 41:27 primus ad Sion dicet ecce adsunt I will give to Jerusalem one that bringeth good tidings.

41:28 For I beheld, and there was no man; even among et Hierusalem euangelistam dabo 41:28 et uidi et non erat

them, and there was no counsellor, that, when I asked of them, could answer a word.

neque ex istis quisquam qui iniret consilium et interrogatus

41:29 Behold, they are all vanity; their works are nothing: responderet uerbum 41:29 ecce omnes iniusti et uana opera their molten images are wind and confusion.

eorum uentus et inane simulacra eorum

42:1 Behold my servant, whom I uphold; mine elect, in 42:1 ecce seruus meus suscipiam eum electus meus con-

whom my soul delighteth; I have put my spirit upon him: he shall bring forth judgment to the Gentiles.

placuit sibi in illo anima mea dedi spiritum meum super eum

42:2 He shall not cry, nor lift up, nor cause his voice to be iudicium gentibus proferet 42:2 non clamabit neque acci-

heard in the street.

42:3 A bruised reed shall he not break, and the smoking piet personam nec audietur foris uox eius 42:3 calamum

flax shall he not quench: he shall bring forth judgment unto truth.

quassatum non conteret et linum fumigans non extinguet in

42:4 He shall not fail nor be discouraged, till he have set ueritate educet iudicium 42:4 non erit tristis neque turbu-

judgment in the earth: and the isles shall wait for his law.

lentus donec ponat in terra iudicium et legem eius insulae

42:5 Thus saith God the LORD, he that created the heav- expectabunt 42:5 haec dicit Dominus Deus creans caelos et

ens, and stretched them out; he that spread forth the earth, and that which cometh out of it; he that giveth breath unto the people upon it, and spirit to them that walk therein: extendens eos firmans terram et quae germinant ex ea dans

flatum populo qui est super eam et spiritum calcantibus eam

42:6 I the LORD have called thee in righteousness, and 42:6 ego Dominus uocaui te in iustitia et adprehendi manum

will hold thine hand, and will keep thee, and give thee for a covenant of the people, for a light of the Gentiles;

tuam et seruaui et dedi te in foedus populi in lucem gentium

42:7 To open the blind eyes, to bring out the prisoners 42:7 ut aperires oculos caecorum et educeres de conclusione

from the prison, and them that sit in darkness out of the prison house.

uinctum de domo carceris sedentes in tenebris 42:8 ego Do-

42:8 I am the LORD: that is my name: and my glory will I not give to another, neither my praise to graven images.

minus hoc est nomen meum gloriam meam alteri non dabo et

42:9 Behold, the former things are come to pass, and new laudem meam sculptilibus 42:9 quae prima fuerant ecce ue-

things do I declare: before they spring forth I tell you of them.

nerunt noua quoque ego adnuntio antequam oriantur audita

42:10 Sing unto the LORD a new song, and his praise from uobis faciam 42:10 cantate Domino canticum nouum laus

the end of the earth, ye that go down to the sea, and all that is therein; the isles, and the inhabitants thereof.

eius ab extremis terrae qui descenditis in mare et plenitudo

42:11 Let the wilderness and the cities thereof lift up their eius insulae et habitatores earum 42:11 subleuetur desertum

voice, the villages that Kedar doth inhabit: let the inhabitants of the rock sing, let them shout from the top of the

mountains. et ciuitates eius in domibus habitabit Cedar laudate habi-42:12 Let them give glory unto the LORD, and declare his tatores Petrae de uertice montium clamabunt 42:12 ponent

praise in the islands.

42:13 The LORD shall go forth as a mighty man, he shall Domino gloriam et laudem eius in insulis nuntiabunt 42:13

stir up jealousy like a man of war: he shall cry, yea, roar; he shall prevail against his enemies.

Dominus sicut fortis egredietur sicut uir proeliator suscitabit zelum uociferabitur et clamabit super inimicos suos con-

42:14 I have long time holden my peace; I have been fortabitur 42:14 tacui semper silui patiens fui sicut pariens

still, and refrained myself: now will I cry like a travailing woman; I will destroy and devour at once.

loquar dissipabo et absorbebo simul 42:15 desertos faciam

42:15 I will make waste mountains and hills, and dry up all their herbs; and I will make the rivers islands, and I will dry up the pools.

montes et colles et omne gramen eorum exsiccabo et ponam

Prophetia Isaiae

flumina in insulas et stagna arefaciam 42:16 et ducam cae- 42:16 And I will bring the blind by a way that they knew

not; I will lead them in paths that they have not known: I will make darkness light before them, and crooked things straight. These things will I do unto them, and not forsake them.

cos in uia quam nesciunt in semitis quas ignorauerunt ambulare eos faciam ponam tenebras coram eis in lucem et praua in recta haec uerba feci eis et non dereliqui eos 42:17 con- 42:17 They shall be turned back, they shall be greatly

ashamed, that trust in graven images, that say to the molten images, Ye are our gods.

uersi sunt retrorsum confundantur confusione qui confidunt in sculptili qui dicunt conflatili uos dii nostri 42:18 surdi 42:18 Hear, ye deaf; and look, ye blind, that ye may see. audite et caeci intuemini ad uidendum 42:19 quis caecus 42:19 Who is blind, but my servant? or deaf, as my mes-

senger that I sent? who is blind as he that is perfect, and blind as the LORD’s servant?

nisi seruus meus et surdus nisi ad quem nuntios meos misi quis caecus nisi qui uenundatus est quis caecus nisi seruus

Domini 42:20 qui uides multa nonne custodies qui apertas 42:20 Seeing many things, but thou observest not; opening

the ears, but he heareth not.

habes aures nonne audies 42:21 et Dominus uoluit ut sanc- 42:21 The LORD is well pleased for his righteousness’

sake; he will magnify the law, and make it honourable.

tificaret eum et magnificaret legem et extolleret 42:22 ipse 42:22 But this is a people robbed and spoiled; they are all

of them snared in holes, and they are hid in prison houses: they are for a prey, and none delivereth; for a spoil, and

autem populus direptus et uastatus laqueus iuuenum omnes none saith, Restore. et in domibus carcerum absconditi sunt facti sunt in rapinam nec est qui eruat in direptionem et non est qui dicat redde 42:23 quis est in uobis qui audiat hoc adtendat et auscultet 42:23 Who among you will give ear to this? who will

hearken and hear for the time to come?

futura 42:24 quis dedit in direptionem Iacob et Israhel ua- 42:24 Who gave Jacob for a spoil, and Israel to the rob-

bers? did not the LORD, he against whom we have sinned? for they would not walk in his ways, neither were

stantibus nonne Dominus ipse cui peccauimus et noluerunt they obedient unto his law. in uiis eius ambulare et non audierunt legem eius 42:25 et 42:25 Therefore he hath poured upon him the fury of his

anger, and the strength of battle: and it hath set him on fire round about, yet he knew not; and it burned him, yet he

effudit super eum indignationem furoris sui et forte bellum laid it not to heart. et conbusit eum in circuitu et non cognouit et succendit eum et non intellexit

43:1 et nunc haec dicit Dominus creans te Iacob et for- 43:1 But now thus saith the LORD that created thee, O

Jacob, and he that formed thee, O Israel, Fear not: for I have redeemed thee, I have called thee by thy name; thou

mans te Israhel noli timere quia redeei te et uocaui nomine art mine. tuo meus es tu 43:2 cum transieris per aquas tecum ero et 43:2 When thou passest through the waters, I will be

with thee; and through the rivers, they shall not overflow thee: when thou walkest through the fire, thou shalt not be

flumina non operient te cum ambulaueris in igne non con- burned; neither shall the flame kindle upon thee. bureris et flamma non ardebit in te 43:3 quia ego Dominus 43:3 For I am the LORD thy God, the Holy One of Israel,

thy Saviour: I gave Egypt for thy ransom, Ethiopia and Seba for thee.

Deus tuus Sanctus Israhel saluator tuus dedi propitiationem

tuam Aegyptum Aethiopiam et Saba pro te 43:4 ex quo ho- 43:4 Since thou wast precious in my sight, thou hast been honourable, and I have loved thee: therefore will I give men for thee, and people for thy life.

norabilis factus es in oculis meis et gloriosus ego dilexi te et dabo homines pro te et populos pro anima tua 43:5 noli 43:5 Fear not: for I am with thee: I will bring thy seed

from the east, and gather thee from the west;

timere quoniam tecum ego sum ab oriente adducam semen tuum et ab occidente congregabo te 43:6 dicam aquiloni da 43:6 I will say to the north, Give up; and to the south, Keep

not back: bring my sons from far, and my daughters from the ends of the earth;

Biblia Sacra Vulgata

et austro noli prohibere adfer filios meos de longinquo et fi-

43:7 Even every one that is called by my name: for I have lias meas ab extremis terrae 43:7 et omnem qui inuocat no-

created him for my glory, I have formed him; yea, I have made him.

men meum in gloriam meam creaui eum et formaui eum et

43:8 Bring forth the blind people that have eyes, and the feci eum 43:8 educ foras populum caecum et oculos habendeaf that have ears. 43:9 Let all the nations be gathered together, and let the tem surdum et aures ei sunt 43:9 omnes gentes congregatae people be assembled: who among them can declare this, and shew us former things? let them bring forth their witnesses, that they may be justified: or let them hear, and say, It is truth.

sunt simul et collectae sunt tribus quis in uobis adnuntiet istud et quae prima sunt audire nos faciat dent testes eorum et

43:10 Ye are my witnesses, saith the LORD, and my ser- iustificentur et audiant et dicant uere 43:10 uos testes mei

vant whom I have chosen: that ye may know and believe me, and understand that I am he: before me there was no

God formed, neither shall there be after me. dicit Dominus et seruus meus quem elegi ut sciatis et cre-

datis mihi et intellegatis quia ego ipse sum ante me non est

43:11 I, even I, am the LORD; and beside me there is no formatus deus et post me non erit 43:11 ego sum ego sum

saviour. 43:12 I have declared, and have saved, and I have shewed, Dominus et non est absque me saluator 43:12 ego adnunwhen there was no strange god among you: therefore ye are my witnesses, saith the LORD, that I am God.

tiaui et saluaui auditum feci et non fuit in uobis alienus uos

43:13 Yea, before the day was I am he; and there is none testes mei dicit Dominus et ego Deus 43:13 et ab initio ego

that can deliver out of my hand: I will work, and who shall let it?

ipse et non est qui de manu mea eruat operabor et quis auer-

43:14 Thus saith the LORD, your redeemer, the Holy One tet illud 43:14 haec dicit Dominus redeeptor uester Sanctus

of Israel; For your sake I have sent to Babylon, and have brought down all their nobles, and the Chaldeans, whose

cry is in the ships. Israhel propter uos emisi Babylonem et detraxi uectes uni-43:15 I am the LORD, your Holy One, the creator of Is- uersos et Chaldeos in nauibus suis gloriantes 43:15 ego Do-

rael, your King.

43:16 Thus saith the LORD, which maketh a way in the minus Sanctus uester creans Israhel Rex uester 43:16 haec

sea, and a path in the mighty waters;

dicit Dominus qui dedit in mari uiam et in aquis torrenti-

43:17 Which bringeth forth the chariot and horse, the army bus semitam 43:17 qui eduxit quadrigam et equum agmen and the power; they shall lie down together, they shall not rise: they are extinct, they are quenched as tow.

et robustum simul obdormierunt nec resurgent contriti sunt

43:18 Remember ye not the former things, neither con- quasi linum et extincti sunt 43:18 ne memineritis priorum

sider the things of old.

43:19 Behold, I will do a new thing; now it shall spring et antiqua ne intueamini 43:19 ecce ego facio noua et nunc

forth; shall ye not know it? I will even make a way in the wilderness, and rivers in the desert.

orientur utique cognoscetis ea ponam in deserto uiam et in

43:20 The beast of the field shall honour me, the dragons inuio flumina 43:20 glorificabit me bestia agri dracones et

and the owls: because I give waters in the wilderness, and rivers in the desert, to give drink to my people, my chosen.

strutiones quia dedi in deserto aquas flumina in inuio ut da-

43:21 This people have I formed for myself; they shall rem potum populo meo electo meo 43:21 populum istum shew forth my praise. 43:22 But thou hast not called upon me, O Jacob; but thou formaui mihi laudem meam narrabit 43:22 non me inuohast been weary of me, O Israel.

43:23 Thou hast not brought me the small cattle of thy casti Iacob nec laborasti in me Israhel 43:23 non obtulisti

burnt offerings; neither hast thou honoured me with thy sacrifices. I have not caused thee to serve with an offering,

nor wearied thee with incense. mihi arietem holocausti tui et uictimis tuis non glorificasti

me non te seruire feci in oblatione nec laborem tibi prae-

43:24 Thou hast bought me no sweet cane with money, bui in ture 43:24 non emisti mihi argento calamum et adipe

neither hast thou filled me with the fat of thy sacrifices: but thou hast made me to serve with thy sins, thou hast wearied me with thine iniquities.

Prophetia Isaiae

uictimarum tuarum non inebriasti me uerumtamen seruire me fecisti in peccatis tuis praebuisti mihi laborem in iniqui-

tatibus tuis 43:25 ego sum ego sum ipse qui deleo iniqui- 43:25 I, even I, am he that blotteth out thy transgressions

for mine own sake, and will not remember thy sins.

tates tuas propter me et peccatorum tuorum non recordabor 43:26 reduc me in memoriam et iudicemur simul narra si quid 43:26 Put me in remembrance: let us plead together: de-

clare thou, that thou mayest be justified.

habes ut iustificeris 43:27 pater tuus primus peccauit et in- 43:27 Thy first father hath sinned, and thy teachers have

transgressed against me.

terpretes tui praeuaricati sunt in me 43:28 et contaminaui 43:28 Therefore I have profaned the princes of the sanc-

tuary, and have given Jacob to the curse, and Israel to reproaches.

principes sanctos dedi ad internicionem Iacob et Israhel in blasphemiam

44:1 et nunc audi Iacob serue meus et Israhel quem elegi 44:1 Yet now hear, O Jacob my servant; and Israel, whom

I have chosen:

44:2 haec dicit Dominus faciens et formans te ab utero auxi- 44:2 Thus saith the LORD that made thee, and formed thee

from the womb, which will help thee; Fear not, O Jacob, my servant; and thou, Jesurun, whom I have chosen.

liator tuus noli timere serue meus Iacob et Rectissime quem elegi 44:3 effundam enim aquas super sitientem et fluenta 44:3 For I will pour water upon him that is thirsty, and

floods upon the dry ground: I will pour my spirit upon thy seed, and my blessing upon thine offspring:

super aridam effundam spiritum meum super semen tuum et benedictionem meam super stirpem tuam 44:4 et germi- 44:4 And they shall spring up as among the grass, as wil-

lows by the water courses.

nabunt inter herbas quasi salices iuxta praeterfluentes aquas 44:5 iste dicet Domini ego sum et ille uocabit in nomine Ia- 44:5 One shall say, I am the LORD’s; and another shall

call himself by the name of Jacob; and another shall subscribe with his hand unto the LORD, and surname himself

cob et hic scribet manu sua Domino et in nomine Israhel by the name of Israel. adsimilabitur 44:6 haec dicit Dominus rex Israhel et rede- 44:6 Thus saith the LORD the King of Israel, and his re-

deemer the LORD of hosts; I am the first, and I am the last; and beside me there is no God.

eptor eius Dominus exercituum ego primus et ego nouissi-

mus et absque me non est deus 44:7 quis similis mei uo- 44:7 And who, as I, shall call, and shall declare it, and set it in order for me, since I appointed the ancient people? and the things that are coming, and shall come, let them

cet et adnuntiet et ordinem exponat mihi ex quo constitui shew unto them. populum antiquum uentura et quae futura sunt adnuntient eis 44:8 nolite timere neque conturbemini ex tunc audire 44:8 Fear ye not, neither be afraid: have not I told thee

from that time, and have declared it? ye are even my witnesses. Is there a God beside me? yea, there is no God; I

te feci et adnuntiaui uos estis testes mei numquid est deus know not any.

absque me et formator quem ego non nouerim 44:9 plastae 44:9 They that make a graven image are all of them vanity; and their delectable things shall not profit; and they are their own witnesses; they see not, nor know; that they may

idoli omnes nihil sunt et amantissima eorum non prodeeunt be ashamed. eis ipsi sunt testes eorum quia non uident neque intellegunt ut confundantur 44:10 quis formauit deum et sculptile con- 44:10 Who hath formed a god, or molten a graven image

that is profitable for nothing?

flauit ad nihil utile 44:11 ecce omnes participes eius con- 44:11 Behold, all his fellows shall be ashamed: and the

workmen, they are of men: let them all be gathered together, let them stand up; yet they shall fear, and they shall

fundentur fabri enim sunt ex hominibus conuenient omnes be ashamed together. stabunt et pauebunt et confundentur simul 44:12 faber fer- 44:12 The smith with the tongs both worketh in the coals,

and fashioneth it with hammers, and worketh it with the strength of his arms: yea, he is hungry, and his strength

rarius lima operatus est in prunis et in malleis formauit il- faileth: he drinketh no water, and is faint.

Biblia Sacra Vulgata

lud et operatus est in brachio fortitudinis suae esuriet et de-

44:13 The carpenter stretcheth out his rule; he marketh it ficiet non bibet aquam et lassescet 44:13 artifex lignarius

out with a line; he fitteth it with planes, and he marketh it out with the compass, and maketh it after the figure of a man, according to the beauty of a man; that it may remain in the house.

extendit normam formauit illud in runcina fecit illud in angularibus et in circino tornauit illud et fecit imaginem uiri

44:14 He heweth him down cedars, and taketh the cypress quasi speciosum hominem habitantem in domo 44:14 sucand the oak, which he strengtheneth for himself among the trees of the forest: he planteth an ash, and the rain doth

nourish it. cidit cedros tulit ilicem et quercum quae steterat inter ligna 44:15 Then shall it be for a man to burn: for he will saltus plantauit pinum quam pluuia nutriuit 44:15 et facta

take thereof, and warm himself; yea, he kindleth it, and baketh bread; yea, he maketh a god, and worshippeth it;

he maketh it a graven image, and falleth down thereto. est hominibus in focum sumpsit ex eis et calefactus est et

succendit et coxit panes de reliquo autem operatus est deum

44:16 He burneth part thereof in the fire; with part thereof et adorauit fecit sculptile et curuatus est ante illud 44:16

he eateth flesh; he roasteth roast, and is satisfied: yea, he warmeth himself, and saith, Aha, I am warm, I have seen

the fire: medium eius conbusit igni et de medio eius carnes comedit

coxit pulmentum et saturatus est et calefactus est et dixit ua

44:17 And the residue thereof he maketh a god, even his calefactus sum uidi focum 44:17 reliquum autem eius deum

graven image: he falleth down unto it, and worshippeth it, and prayeth unto it, and saith, Deliver me; for thou art my

god. fecit sculptile sibi curuatur ante illud et adorat illud et obse-44:18 They have not known nor understood: for he hath crat dicens libera me quia deus meus es tu 44:18 nescierunt

shut their eyes, that they cannot see; and their hearts, that they cannot understand.

neque intellexerunt lutati enim sunt ne uideant oculi eorum

44:19 And none considereth in his heart, neither is there et ne intellegant corde suo 44:19 non recogitant in mente knowledge nor understanding to say, I have burned part of it in the fire; yea, also I have baked bread upon the coals thereof; I have roasted flesh, and eaten it: and shall I make the residue thereof an abomination? shall I fall down to the stock of a tree?

sua neque cognoscunt neque sentiunt ut dicant medietatem eius conbusi igne et coxi super carbones eius panes coxi carnes et comedi et de reliquo eius idolum faciam ante truncum

44:20 He feedeth on ashes: a deceived heart hath turned ligni procidam 44:20 pars eius cinis est cor insipiens ad-

him aside, that he cannot deliver his soul, nor say, Is there not a lie in my right hand?

orauit illud et non liberabit animam suam neque dicet forte

44:21 Remember these, O Jacob and Israel; for thou art mendacium est in dextera mea 44:21 memento horum Iamy servant: I have formed thee; thou art my servant: O Israel, thou shalt not be forgotten of me.

cob et Israhel quoniam seruus meus es tu formaui te seruus

44:22 I have blotted out, as a thick cloud, thy transgres- meus es tu Israhel non obliuisceris mei 44:22 deleui ut nu-

sions, and, as a cloud, thy sins: return unto me; for I have redeemed thee.

bem iniquitates tuas et quasi nebulam peccata tua reuertere

44:23 Sing, O ye heavens; for the LORD hath done it: ad me quoniam redeei te 44:23 laudate caeli quoniam fecit

shout, ye lower parts of the earth: break forth into singing, ye mountains, O forest, and every tree therein: for the LORD hath redeemed Jacob, and glorified himself in Israel.

Dominus iubilate extrema terrae resonate montes laudationem saltus et omne lignum eius quoniam redeeit Dominus

44:24 Thus saith the LORD, thy redeemer, and he that Iacob et Israhel gloriabitur 44:24 haec dicit Dominus red-

formed thee from the womb, I am the LORD that maketh all things; that stretcheth forth the heavens alone; that

spreadeth abroad the earth by myself; eeptor tuus et formator tuus ex utero ego sum Dominus fa-

ciens omnia extendens caelos solus stabiliens terram et nul-

44:25 That frustrateth the tokens of the liars, and maketh lus mecum 44:25 irrita faciens signa diuinorum et ariolos in

diviners mad; that turneth wise men backward, and maketh their knowledge foolish;

Prophetia Isaiae

furorem uertens conuertens sapientes retrorsum et scientiam eorum stultam faciens 44:26 suscitans uerbum serui sui et 44:26 That confirmeth the word of his servant, and per-

formeth the counsel of his messengers; that saith to Jerusalem, Thou shalt be inhabited; and to the cities of Judah, Ye shall be built, and I will raise up the decayed places thereof:

consilium nuntiorum suorum conplens qui dico Hierusalem habitaberis et ciuitatibus Iuda aedificabimini et deserta eius suscitabo 44:27 qui dico profundo desolare et flumina tua 44:27 That saith to the deep, Be dry, and I will dry up thy

rivers:

arefaciam 44:28 qui dico Cyro pastor meus es et omnem 44:28 That saith of Cyrus, He is my shepherd, and shall perform all my pleasure: even saying to Jerusalem, Thou shalt be built; and to the temple, Thy foundation shall be

uoluntatem meam conplebis qui dico Hierusalem aedifica- laid. beris et templo fundaberis

45:1 haec dicit Dominus christo meo Cyro cuius adpre- 45:1 Thus saith the LORD to his anointed, to Cyrus,

whose right hand I have holden, to subdue nations before him; and I will loose the loins of kings, to open before him

hendi dexteram ut subiciam ante faciem eius gentes et dorsa the two leaved gates; and the gates shall not be shut; regum uertam et aperiam coram eo ianuas et portae non cludentur 45:2 ego ante te ibo et gloriosos terrae humiliabo 45:2 I will go before thee, and make the crooked places

straight: I will break in pieces the gates of brass, and cut in sunder the bars of iron:

portas aereas conteram et uectes ferreos confringam 45:3

45:3 And I will give thee the treasures of darkness, and hidden riches of secret places, that thou mayest know that I, the LORD, which call thee by thy name, am the God of Israel.

et dabo tibi thesauros absconditos et arcana secretorum ut scias quia ego Dominus qui uoco nomen tuum Deus Israhel

45:4 propter seruum meum Iacob et Israhel electum meum 45:4 For Jacob my servant’s sake, and Israel mine elect, I have even called thee by thy name: I have surnamed thee, though thou hast not known me.

et uocaui te in nomine tuo adsimilaui te et non cognouisti me 45:5 ego Dominus et non est amplius extra me non est 45:5 I am the LORD, and there is none else, there is no

God beside me: I girded thee, though thou hast not known me:

deus accinxi te et non cognouisti me 45:6 ut sciant hii qui

45:6 That they may know from the rising of the sun, and from the west, that there is none beside me. I am the LORD, and there is none else.

ab ortu solis et qui ab occidente quoniam absque me non est ego Dominus et non est alter 45:7 formans lucem et creans 45:7 I form the light, and create darkness: I make peace,

and create evil: I the LORD do all these things.

tenebras faciens pacem et creans malum ego Dominus faciens omnia haec 45:8 rorate caeli desuper et nubes pluant 45:8 Drop down, ye heavens, from above, and let the skies

pour down righteousness: let the earth open, and let them bring forth salvation, and let righteousness spring up to-

iustum aperiatur terra et germinet saluatorem et iustitia oria- gether; I the LORD have created it. tur simul ego Dominus creaui eum 45:9 uae qui contradicit 45:9 Woe unto him that striveth with his Maker! Let the

potsherd strive with the potsherds of the earth. Shall the clay say to him that fashioneth it, What makest thou? or

fictori suo testa de samiis terrae numquid dicet lutum figulo thy work, He hath no hands? suo quid facis et opus tuum absque manibus est 45:10 uae 45:10 Woe unto him that saith unto his father, What

begettest thou? or to the woman, What hast thou brought forth?

qui dicit patri quid generas et mulieri quid parturis 45:11

45:11 Thus saith the LORD, the Holy One of Israel, and his Maker, Ask me of things to come concerning my sons, and concerning the work of my hands command ye me.

haec dicit Dominus Sanctus Israhel plastes eius uentura interrogate me super filios meos et super opus manuum mearum mandastis mihi 45:12 ego feci terram et hominem su- 45:12 I have made the earth, and created man upon it: I,

even my hands, have stretched out the heavens, and all their host have I commanded.

per eam creaui ego manus meae tetenderunt caelos et omni militiae eorum mandaui 45:13 ego suscitaui eum ad iusti- 45:13 I have raised him up in righteousness, and I will

direct all his ways: he shall build my city, and he shall let go my captives, not for price nor reward, saith the LORD of hosts.

Biblia Sacra Vulgata

tiam et omnes uias eius dirigam ipse aedificabit ciuitatem meam et captiuitatem meam dimittet non in pretio neque in

45:14 Thus saith the LORD, The labour of Egypt, and muneribus dicit Dominus Deus exercituum 45:14 haec dicit

merchandise of Ethiopia and of the Sabeans, men of stature, shall come over unto thee, and they shall be thine: they shall come after thee; in chains they shall come over, and they shall fall down unto thee, they shall make supplication unto thee, saying, Surely God is in thee; and there is none else, there is no God.

Dominus labor Aegypti et negotiatio Aethiopiae et Sabaim uiri sublimes ad te transibunt et tui erunt post te ambulabunt uincti manicis pergent et te adorabunt teque deprecabuntur

45:15 Verily thou art a God that hidest thyself, O God of tantum in te est Deus et non est absque te deus 45:15 uere Israel, the Saviour.

45:16 They shall be ashamed, and also confounded, all of tu es Deus absconditus Deus Israhel saluator 45:16 confusi

them: they shall go to confusion together that are makers of idols.

sunt et erubuerunt omnes simul abierunt in confusione fabri-

45:17 But Israel shall be saved in the LORD with an ever- catores errorum 45:17 Israhel saluatus est in Domino salute

lasting salvation: ye shall not be ashamed nor confounded world without end.

aeterna non confundemini et non erubescetis usque in sae-

45:18 For thus saith the LORD that created the heavens; culum saeculi 45:18 quia haec dicit Dominus creans caelos

God himself that formed the earth and made it; he hath established it, he created it not in vain, he formed it to be

inhabited: I am the LORD; and there is none else. ipse Deus formans terram et faciens eam ipse plastes eius

non in uanum creauit eam ut habitetur formauit eam ego

45:19 I have not spoken in secret, in a dark place of the Dominus et non est alius 45:19 non in abscondito locutus

earth: I said not unto the seed of Jacob, Seek ye me in vain: I the LORD speak righteousness, I declare things

that are right. sum in loco terrae tenebroso non dixi semini Iacob frustra

quaerite me ego Dominus loquens iustitiam adnuntians recta

45:20 Assemble yourselves and come; draw near together, 45:20 congregamini et uenite et accedite simul qui saluati es-

ye that are escaped of the nations: they have no knowledge that set up the wood of their graven image, and pray unto

a god that cannot save. tis ex gentibus nescierunt qui leuant lignum sculpturae suae 45:21 Tell ye, and bring them near; yea, let them take et rogant deum non saluantem 45:21 adnuntiate et uenite et

counsel together: who hath declared this from ancient time? who hath told it from that time? have not I the LORD? and there is no God else beside me; a just God and a Saviour; there is none beside me.

consiliamini simul quis auditum fecit hoc ab initio ex tunc praedixit illud numquid non ego Dominus et non est ultra Deus absque me Deus iustus et saluans non est praeter me

45:22 Look unto me, and be ye saved, all the ends of the 45:22 conuertimini ad me et salui eritis omnes fines terrae

earth: for I am God, and there is none else.

45:23 I have sworn by myself, the word is gone out of my quia ego Deus et non est alius 45:23 in memet ipso iuraui

mouth in righteousness, and shall not return, That unto me every knee shall bow, every tongue shall swear.

egredietur de ore meo iustitiae uerbum et non reuertetur quia

45:24 Surely, shall one say, in the LORD have I righteous- mihi curuabunt omnia genu et iurabit omnis lingua 45:24 ness and strength: even to him shall men come; and all that are incensed against him shall be ashamed.

ergo in Domino dicet meae sunt iustitiae et imperium ad

45:25 In the LORD shall all the seed of Israel be justified, eum uenient et confundentur omnes qui repugnant ei 45:25 and shall glory.

in Domino iustificabitur et laudabitur omne semen Israhel

46:1 Bel boweth down, Nebo stoopeth, their idols were 46:1 conflatus est Bel contritus est Nabo facta sunt simu-

upon the beasts, and upon the cattle: your carriages were heavy loaden; they are a burden to the weary beast.

lacra eorum bestiis et iumentis onera uestra graui pondere

46:2 They stoop, they bow down together; they could not usque ad lassitudinem 46:2 contabuerunt et contrita sunt si-

deliver the burden, but themselves are gone into captivity.

Prophetia Isaiae

mul non potuerunt saluare portantem et anima eorum in captiuitatem ibit 46:3 audite me domus Iacob et omne resi- 46:3 Hearken unto me, O house of Jacob, and all the rem-

nant of the house of Israel, which are borne by me from the belly, which are carried from the womb:

duum domus Israhel qui portamini a meo utero qui gestamini a mea uulua 46:4 usque ad senectam ego ipse et usque 46:4 And even to your old age I am he; and even to hoar

hairs will I carry you: I have made, and I will bear; even I will carry, and will deliver you.

ad canos ego portabo ego feci et ego feram et ego portabo et saluabo 46:5 cui adsimilastis me et adaequastis et con- 46:5 To whom will ye liken me, and make me equal, and

compare me, that we may be like?

parastis me et fecistis similem 46:6 qui confertis aurum de 46:6 They lavish gold out of the bag, and weigh silver in

the balance, and hire a goldsmith; and he maketh it a god: they fall down, yea, they worship.

sacculo et argentum statera ponderatis conducentes aurificem ut faciat deum et procidunt et adorant 46:7 portant il- 46:7 They bear him upon the shoulder, they carry him, and

set him in his place, and he standeth; from his place shall he not remove: yea, one shall cry unto him, yet can he not

lud in umeris gestantes et ponentes in loco suo et stabit ac answer, nor save him out of his trouble. de loco suo non mouebitur sed et cum clamauerint ad eum non audiet de tribulatione non saluabit eos 46:8 memen- 46:8 Remember this, and shew yourselves men: bring it

again to mind, O ye transgressors.

tote istud et fundamini rediie praeuaricatores ad cor 46:9 46:9 Remember the former things of old: for I am God,

and there is none else; I am God, and there is none like me,

recordamini prioris saeculi quoniam ego sum Deus et non est ultra Deus nec est similis mei 46:10 adnuntians ab exor- 46:10 Declaring the end from the beginning, and from an-

cient times the things that are not yet done, saying, My counsel shall stand, and I will do all my pleasure:

dio nouissimum et ab initio quae necdum facta sunt dicens consilium meum stabit et omnis uoluntas mea fiet 46:11 uo- 46:11 Calling a ravenous bird from the east, the man that

executeth my counsel from a far country: yea, I have spoken it, I will also bring it to pass; I have purposed it, I will

cans ab oriente auem et de terra longinqua uirum uoluntatis also do it. meae et locutus sum et adducam illud creaui et faciam illud 46:12 audite me duro corde qui longe estis a iustitia 46:13 46:12 Hearken unto me, ye stouthearted, that are far from

righteousness:

46:13 I bring near my righteousness; it shall not be far off, and my salvation shall not tarry: and I will place salvation in Zion for Israel my glory.

prope feci iustitiam meam non elongabitur et salus mea non morabitur dabo in Sion salutem et Israheli gloriam meam

47:1 descende sede in puluerem uirgo filia Babylon sede in 47:1 Come down, and sit in the dust, O virgin daughter of

Babylon, sit on the ground: there is no throne, O daughter of the Chaldeans: for thou shalt no more be called tender

terra non est solium filiae Chaldeorum quia ultra non uoca- and delicate. beris mollis et tenera 47:2 tolle molam et mole farinam de- 47:2 Take the millstones, and grind meal: uncover thy

locks, make bare the leg, uncover the thigh, pass over the rivers.

nuda turpitudinem tuam discoperi umerum reuela crus transi flumina 47:3 reuelabitur ignominia tua et uidebitur obprob- 47:3 Thy nakedness shall be uncovered, yea, thy shame

shall be seen: I will take vengeance, and I will not meet thee as a man.

rium tuum ultionem capiam et non resistet mihi homo 47:4

47:4 As for our redeemer, the LORD of hosts is his name, the Holy One of Israel.

redeeptor noster Dominus exercituum nomen illius Sanctus Israhel 47:5 sede tace et intra in tenebras filia Chaldeorum 47:5 Sit thou silent, and get thee into darkness, O daughter

of the Chaldeans: for thou shalt no more be called, The lady of kingdoms.

quia non uocaberis ultra domina regnorum 47:6 iratus sum

47:6 I was wroth with my people, I have polluted mine inheritance, and given them into thine hand: thou didst shew them no mercy; upon the ancient hast thou very heavily laid thy yoke.

super populum meum contaminaui hereditatem meam et dedi eos in manu tua non posuisti eis misericordias super se-

Biblia Sacra Vulgata

47:7 And thou saidst, I shall be a lady for ever: so that nem adgrauasti iugum tuum ualde 47:7 et dixisti in sempi-

thou didst not lay these things to thy heart, neither didst remember the latter end of it.

ternum ero domina non posuisti haec super cor tuum neque

47:8 Therefore hear now this, thou that art given to plea- recordata es nouissimi tui 47:8 et nunc audi haec delicata et

sures, that dwellest carelessly, that sayest in thine heart, I am, and none else beside me; I shall not sit as a widow,

neither shall I know the loss of children: habitans confidenter quae dicis in corde tuo ego sum et non

est praeter me amplius non sedebo uidua et ignorabo sterili-

47:9 But these two things shall come to thee in a moment tatem 47:9 uenient tibi duo haec subito in die una sterilitas

in one day, the loss of children, and widowhood: they shall come upon thee in their perfection for the multitude of thy sorceries, and for the great abundance of thine enchantments.

et uiduitas uniuersa uenerunt super te propter multitudinem maleficiorum tuorum et propter duritiam incantatorum tuo-

47:10 For thou hast trusted in thy wickedness: thou hast rum uehementem 47:10 et fiduciam habuisti in malitia tua said, None seeth me. Thy wisdom and thy knowledge, it hath perverted thee; and thou hast said in thine heart, I am,

and none else beside me. et dixisti non est qui uideat me sapientia tua et scientia tua

haec decepit te et dixisti in corde tuo ego sum et praeter me

47:11 Therefore shall evil come upon thee; thou shalt not non est altera 47:11 ueniet super te malum et nescies ortum

know from whence it riseth: and mischief shall fall upon thee; thou shalt not be able to put it off: and desolation

shall come upon thee suddenly, which thou shalt not know. eius et inruet super te calamitas quam non poteris expiare 47:12 Stand now with thine enchantments, and with the ueniet super te repente miseria quam nescies 47:12 sta cum

multitude of thy sorceries, wherein thou hast laboured from thy youth; if so be thou shalt be able to profit, if

so be thou mayest prevail. incantatoribus tuis et cum multitudine maleficiorum tuorum

in quibus laborasti ab adulescentia tua si forte quid prosit

47:13 Thou art wearied in the multitude of thy counsels. tibi aut si possis fieri fortior 47:13 defecisti in multitudine

Let now the astrologers, the stargazers, the monthly prognosticators, stand up, and save thee from these things that

shall come upon thee. consiliorum tuorum stent et saluent te augures caeli qui con-

templabantur sidera et supputabant menses ut ex eis adnun-

47:14 Behold, they shall be as stubble; the fire shall burn tiarent uentura tibi 47:14 ecce facti sunt quasi stipula ignis

them; they shall not deliver themselves from the power of the flame: there shall not be a coal to warm at, nor fire to

sit before it. conbusit eos non liberabunt animam suam de manu flam-

mae non sunt prunae quibus calefiant nec focus ut sedeant

47:15 Thus shall they be unto thee with whom thou hast ad eum 47:15 sic facta sunt tibi in quibuscumque laboraue-

laboured, even thy merchants, from thy youth: they shall wander every one to his quarter; none shall save thee.

ras negotiatores tui ab adulescentia tua unusquisque in uia sua errauerunt non est qui saluet te

48:1 Hear ye this, O house of Jacob, which are called by 48:1 audite hoc domus Iacob qui uocamini nomine Israhel

the name of Israel, and are come forth out of the waters of Judah, which swear by the name of the LORD, and make mention of the God of Israel, but not in truth, nor in righteousness.

et de aquis Iuda existis qui iuratis in nomine Domini et Dei

48:2 For they call themselves of the holy city, and stay Israhel recordamini non in ueritate neque in iustitia 48:2 de

themselves upon the God of Israel; The LORD of hosts is his name.

ciuitate enim sancta uocati sunt et super Deum Israhel con-

48:3 I have declared the former things from the beginning; stabiliti sunt Dominus exercituum nomen eius 48:3 priora and they went forth out of my mouth, and I shewed them; I did them suddenly, and they came to pass.

ex tunc adnuntiaui et ex ore meo exierunt et audita feci ea

48:4 Because I knew that thou art obstinate, and thy neck repente operatus sum et uenerunt 48:4 sciui enim quia du-

is an iron sinew, and thy brow brass;

48:5 I have even from the beginning declared it to thee; rus es tu et neruus ferreus ceruix tua et frons tua aerea 48:5

before it came to pass I shewed it thee: lest thou shouldest say, Mine idol hath done them, and my graven image, and my molten image, hath commanded them.

Prophetia Isaiae

praedixi tibi ex tunc antequam uenirent indicaui tibi ne forte diceres idola mea fecerunt haec et sculptilia mea et conflatilia mandauerunt ista 48:6 quae audisti uide omnia uos au- 48:6 Thou hast heard, see all this; and will not ye declare

it? I have shewed thee new things from this time, even hidden things, and thou didst not know them.

tem non adnuntiastis audita feci tibi noua ex nunc et conseruata quae nescis 48:7 nunc creata sunt et non ex tunc et 48:7 They are created now, and not from the beginning;

even before the day when thou heardest them not; lest thou shouldest say, Behold, I knew them.

ante diem et non audisti ea ne forte dicas ecce cognoui ea 48:8 neque audisti neque cognouisti neque ex tunc aperta est 48:8 Yea, thou heardest not; yea, thou knewest not; yea,

from that time that thine ear was not opened: for I knew that thou wouldest deal very treacherously, and wast called

auris tua scio enim quia praeuaricans praeuaricabis et trans- a transgressor from the womb. gressorem ex uentre uocaui te 48:9 propter nomen meum 48:9 For my name’s sake will I defer mine anger, and for

my praise will I refrain for thee, that I cut thee not off.

longe faciam furorem meum et laude mea infrenabo te ne intereas 48:10 ecce excoxi te sed non quasi argentum elegi 48:10 Behold, I have refined thee, but not with silver; I

have chosen thee in the furnace of affliction.

te in camino paupertatis 48:11 propter me propter me fa- 48:11 For mine own sake, even for mine own sake, will I

do it: for how should my name be polluted? and I will not give my glory unto another.

ciam ut non blasphemer et gloriam meam alteri non dabo 48:12 audi me Iacob et Israhel quem ego uoco ego ipse ego 48:12 Hearken unto me, O Jacob and Israel, my called; I

am he; I am the first, I also am the last.

primus et ego nouissimus 48:13 manus quoque mea funda- 48:13 Mine hand also hath laid the foundation of the earth, and my right hand hath spanned the heavens: when I call unto them, they stand up together.

uit terram et dextera mea mensa est caelos ego uocabo eos

et stabunt simul 48:14 congregamini omnes uos et audite 48:14 All ye, assemble yourselves, and hear; which among them hath declared these things? The LORD hath loved him: he will do his pleasure on Babylon, and his arm shall

quis de eis adnuntiauit haec Dominus dilexit eum faciet uo- be on the Chaldeans. luntatem suam in Babylone et brachium suum in Chaldeis

48:15 ego ego locutus sum et uocaui eum adduxi eum et di- 48:15 I, even I, have spoken; yea, I have called him: I have

brought him, and he shall make his way prosperous.

recta est uia eius 48:16 accedite ad me et audite hoc non 48:16 Come ye near unto me, hear ye this; I have not spo-

ken in secret from the beginning; from the time that it was, there am I: and now the Lord GOD, and his Spirit, hath

a principio in abscondito locutus sum ex tempore antequam sent me. fieret ibi eram et nunc Dominus Deus misit me et spiritus eius 48:17 haec dicit Dominus redeeptor tuus Sanctus Is- 48:17 Thus saith the LORD, thy Redeemer, the Holy One

of Israel; I am the LORD thy God which teacheth thee to profit, which leadeth thee by the way that thou shouldest

rahel ego Dominus Deus tuus docens te utilia gubernans te go. in uia qua ambulas 48:18 utinam adtendisses mandata mea 48:18 O that thou hadst hearkened to my commandments!

then had thy peace been as a river, and thy righteousness as the waves of the sea:

facta fuisset sicut flumen pax tua et iustitia tua sicut gurgites maris 48:19 et fuisset quasi harena semen tuum et stirps 48:19 Thy seed also had been as the sand, and the off-

spring of thy bowels like the gravel thereof; his name should not have been cut off nor destroyed from before

uteri tui ut lapilli eius non interisset et non fuisset adtritum me. nomen eius a facie mea 48:20 egredimini de Babylone fu- 48:20 Go ye forth of Babylon, flee ye from the Chaldeans,

with a voice of singing declare ye, tell this, utter it even to the end of the earth; say ye, The LORD hath redeemed his

gite a Chaldeis in uoce exultationis adnuntiate auditum facite servant Jacob. hoc efferte illud usque ad extrema terrae dicite redeeit Dominus seruum suum Iacob 48:21 non sitierunt in deserto cum 48:21 And they thirsted not when he led them through the

deserts: he caused the waters to flow out of the rock for them: he clave the rock also, and the waters gushed out.

Biblia Sacra Vulgata

educeret eos aquam de petra produxit eis et scidit petram et

48:22 There is no peace, saith the LORD, unto the wicked. fluxerunt aquae 48:22 non est pax dicit Dominus impiis

49:1 Listen, O isles, unto me; and hearken, ye people, 49:1 audite insulae et adtendite populi de longe Dominus

from far; The LORD hath called me from the womb; from the bowels of my mother hath he made mention of my

name. ab utero uocauit me de uentre matris meae recordatus est no-49:2 And he hath made my mouth like a sharp sword; in minis mei 49:2 et posuit os meum quasi gladium acutum in

the shadow of his hand hath he hid me, and made me a polished shaft; in his quiver hath he hid me;

umbra manus suae protexit me et posuit me sicut sagittam

49:3 And said unto me, Thou art my servant, O Israel, in electam in faretra sua abscondit me 49:3 et dixit mihi ser-

whom I will be glorified.

49:4 Then I said, I have laboured in vain, I have spent my uus meus es tu Israhel quia in te gloriabor 49:4 et ego dixi

strength for nought, and in vain: yet surely my judgment is with the LORD, and my work with my God.

in uacuum laboraui sine causa et uane fortitudinem meam consumpsi ergo iudicium meum cum Domino et opus meum

49:5 And now, saith the LORD that formed me from the cum Deo meo 49:5 et nunc dicit Dominus formans me ex womb to be his servant, to bring Jacob again to him, Though Israel be not gathered, yet shall I be glorious in

the eyes of the LORD, and my God shall be my strength. utero seruum sibi ut reducam Iacob ad eum et Israhel non

congregabitur et glorificatus sum in oculis Domini et Deus

49:6 And he said, It is a light thing that thou shouldest be meus factus est fortitudo mea 49:6 et dixit parum est ut sis

my servant to raise up the tribes of Jacob, and to restore the preserved of Israel: I will also give thee for a light to the Gentiles, that thou mayest be my salvation unto the end of the earth.

mihi seruus ad suscitandas tribus Iacob et feces Israhel conuertendas dedi te in lucem gentium ut sis salus mea usque ad

49:7 Thus saith the LORD, the Redeemer of Israel, and extremum terrae 49:7 haec dicit Dominus redeeptor Isra-

his Holy One, to him whom man despiseth, to him whom the nation abhorreth, to a servant of rulers, Kings shall see and arise, princes also shall worship, because of the LORD that is faithful, and the Holy One of Israel, and he shall choose thee.

hel Sanctus eius ad contemptibilem animam ad abominatam gentem ad seruum dominorum reges uidebunt et consurgent principes et adorabunt propter Dominum quia fidelis est et

49:8 Thus saith the LORD, In an acceptable time have I Sanctum Israhel qui elegit te 49:8 haec dicit Dominus in

heard thee, and in a day of salvation have I helped thee: and I will preserve thee, and give thee for a covenant of the people, to establish the earth, to cause to inherit the desolate heritages;

tempore placito exaudiui te et in die salutis auxiliatus sum tui et seruaui te et dedi te in foedus populi ut suscitares ter-

49:9 That thou mayest say to the prisoners, Go forth; to ram et possideres hereditates dissipatas 49:9 ut diceres his

them that are in darkness, Shew yourselves. They shall feed in the ways, and their pastures shall be in all high

places. qui uincti sunt exite et his qui in tenebris reuelamini super 49:10 They shall not hunger nor thirst; neither shall the uias pascentur et in omnibus planis pascua eorum 49:10 non

heat nor sun smite them: for he that hath mercy on them shall lead them, even by the springs of water shall he guide

them. esurient neque sitient et non percutiet eos aestus et sol quia

miserator eorum reget eos et ad fontes aquarum portabit eos

49:11 And I will make all my mountains a way, and my 49:11 et ponam omnes montes meos in uiam et semitae meae

highways shall be exalted.

49:12 Behold, these shall come from far: and, lo, these exaltabuntur 49:12 ecce isti de longe uenient et ecce illi ab

from the north and from the west; and these from the land of Sinim.

aquilone et mari et isti de terra australi 49:13 laudate caeli

49:13 Sing, O heavens; and be joyful, O earth; and break forth into singing, O mountains: for the LORD hath comforted his people, and will have mercy upon his afflicted.

et exulta terra iubilate montes laudem quia consolatus est Dominus populum suum et pauperum suorum miserebitur

Prophetia Isaiae

49:14 et dixit Sion dereliquit me Dominus et Dominus obli- 49:14 But Zion said, The LORD hath forsaken me, and

my Lord hath forgotten me.

tus est mei 49:15 numquid obliuisci potest mulier infantem 49:15 Can a woman forget her sucking child, that she

should not have compassion on the son of her womb? yea, they may forget, yet will I not forget thee.

suum ut non misereatur filio uteri sui et si illa oblita fuerit ego tamen non obliuiscar tui 49:16 ecce in manibus meis 49:16 Behold, I have graven thee upon the palms of my

hands; thy walls are continually before me.

descripsi te muri tui coram oculis meis semper 49:17 uene- 49:17 Thy children shall make haste; thy destroyers and

they that made thee waste shall go forth of thee.

runt structores tui destruentes te et dissipantes a te exibunt 49:18 leua in circuitu oculos tuos et uide omnes isti congre- 49:18 Lift up thine eyes round about, and behold: all these

gather themselves together, and come to thee. As I live, saith the LORD, thou shalt surely clothe thee with them all, as with an ornament, and bind them on thee, as a bride doeth.

gati sunt uenerunt tibi uiuo ego dicit Dominus quia omnibus his uelut ornamento uestieris et circumdabis tibi eos quasi sponsa 49:19 quia deserta tua et solitudines tuae et terra rui- 49:19 For thy waste and thy desolate places, and the land

of thy destruction, shall even now be too narrow by reason of the inhabitants, and they that swallowed thee up shall

nae tuae nunc angusta erunt prae habitatoribus et longe fu- be far away. gabuntur qui absorbebant te 49:20 adhuc dicent in auribus 49:20 The children which thou shalt have, after thou hast

lost the other, shall say again in thine ears, The place is too strait for me: give place to me that I may dwell.

tuis filii sterilitatis tuae angustus mihi est locus fac spatium mihi ut habitem 49:21 et dices in corde tuo quis genuit mihi 49:21 Then shalt thou say in thine heart, Who hath begot-

ten me these, seeing I have lost my children, and am desolate, a captive, and removing to and fro? and who hath brought up these? Behold, I was left alone; these, where had they been?

istos ego sterilis et non pariens transmigrata et captiua et istos quis enutriuit ego destituta et sola et isti ubi hic erant

49:22 haec dicit Dominus Deus ecce leuo ad gentes manum 49:22 Thus saith the Lord GOD, Behold, I will lift up mine hand to the Gentiles, and set up my standard to the people: and they shall bring thy sons in their arms, and thy daugh-

meam et ad populos exaltabo signum meum et adferent filios ters shall be carried upon their shoulders. tuos in ulnis et filias tuas super umeros portabunt 49:23 et 49:23 And kings shall be thy nursing fathers, and their

queens thy nursing mothers: they shall bow down to thee with their face toward the earth, and lick up the dust of thy feet; and thou shalt know that I am the LORD: for they shall not be ashamed that wait for me.

erunt reges nutricii tui et reginae nutrices tuae uultu in terra dimisso adorabunt te et puluerem pedum tuorum lingent et scies quia ego Dominus super quo non confundentur qui expectant eum 49:24 numquid tolletur a forte praeda aut quod 49:24 Shall the prey be taken from the mighty, or the law-

ful captive delivered?

captum fuerit a robusto saluum esse poterit 49:25 quia haec 49:25 But thus saith the LORD, Even the captives of the

mighty shall be taken away, and the prey of the terrible shall be delivered: for I will contend with him that con-

dicit Dominus equidem et captiuitas a forte tolletur et quod tendeth with thee, and I will save thy children. ablatum fuerit a robusto saluabitur eos uero qui iudicauerunt te ego iudicabo et filios tuos ego saluabo 49:26 et cibabo 49:26 And I will feed them that oppress thee with their

own flesh; and they shall be drunken with their own blood, as with sweet wine: and all flesh shall know that I the LORD am thy Saviour and thy Redeemer, the mighty One of Jacob.

hostes tuos carnibus suis et quasi musto sanguine suo inebriabuntur et sciet omnis caro quia ego Dominus saluans te et redeeptor tuus Fortis Iacob

50:1 haec dicit Dominus quis est hic liber repudii matris 50:1 Thus saith the LORD, Where is the bill of your

mother’s divorcement, whom I have put away? or which of my creditors is it to whom I have sold you? Behold, for your iniquities have ye sold yourselves, and for your transgressions is your mother put away.

uestrae quo dimisi eam aut quis est creditor meus cui uendidi uos ecce in iniquitatibus uestris uenditi estis et in scele-

Biblia Sacra Vulgata

50:2 Wherefore, when I came, was there no man? when I ribus uestris dimisi matrem uestram 50:2 quia ueni et non

called, was there none to answer? Is my hand shortened at all, that it cannot redeem? or have I no power to deliver? behold, at my rebuke I dry up the sea, I make the rivers a wilderness: their fish stinketh, because there is no water, and dieth for thirst.

erat uir uocaui et non erat qui audiret numquid adbreuiata et paruula facta est manus mea ut non possim rediiere aut non est in me uirtus ad liberandum ecce in increpatione mea desertum faciam mare ponam flumina in siccum conputrescent

50:3 I clothe the heavens with blackness, and I make sack- pisces sine aqua et morientur in siti 50:3 induam caelos te-

cloth their covering.

50:4 The Lord GOD hath given me the tongue of the nebris et saccum ponam operimentum eorum 50:4 Domi-

learned, that I should know how to speak a word in season to him that is weary: he wakeneth morning by morning,

he wakeneth mine ear to hear as the learned. nus dedit mihi linguam eruditam ut sciam sustentare eum

qui lassus est uerbo erigit mane mane erigit mihi aurem ut

50:5 The Lord GOD hath opened mine ear, and I was not audiam quasi magistrum 50:5 Dominus Deus aperuit mihi

rebellious, neither turned away back.

50:6 I gave my back to the smiters, and my cheeks to them aurem ego autem non contradico retrorsum non abii 50:6

that plucked off the hair: I hid not my face from shame and spitting.

corpus meum dedi percutientibus et genas meas uellentibus faciem meam non auerti ab increpantibus et conspuentibus

50:7 For the Lord GOD will help me; therefore shall I not 50:7 Dominus Deus auxiliator meus ideo non sum confube confounded: therefore have I set my face like a flint, and I know that I shall not be ashamed.

sus ideo posui faciem meam ut petram durissimam et scio

50:8 He is near that justifieth me; who will contend with quoniam non confundar 50:8 iuxta est qui iustificat me quis

me? let us stand together: who is mine adversary? let him come near to me.

contradicet mihi stemus simul quis est aduersarius meus ac-

50:9 Behold, the Lord GOD will help me; who is he that cedat ad me 50:9 ecce Dominus Deus auxiliator meus quis

shall condemn me? lo, they all shall wax old as a garment; the moth shall eat them up.

est qui condemnet me ecce omnes quasi uestimentum con-

50:10 Who is among you that feareth the LORD, that terentur tinea comedet eos 50:10 quis ex uobis timens Do-

obeyeth the voice of his servant, that walketh in darkness, and hath no light? let him trust in the name of the LORD,

and stay upon his God. minum audiens uocem serui sui qui ambulauit in tenebris et

non est lumen ei speret in nomine Domini et innitatur su-

50:11 Behold, all ye that kindle a fire, that compass your- per Deum suum 50:11 ecce omnes uos accendentes ignem selves about with sparks: walk in the light of your fire, and in the sparks that ye have kindled. This shall ye have of

mine hand; ye shall lie down in sorrow. accincti flammis ambulate in lumine ignis uestri et in flammis quas succendistis de manu mea factum est hoc uobis in doloribus dormietis

51:1 Hearken to me, ye that follow after righteousness, ye 51:1 audite me qui sequimini quod iustum est et quaeritis that seek the LORD: look unto the rock whence ye are hewn, and to the hole of the pit whence ye are digged.

Dominum adtendite ad petram unde excisi estis et ad cauer-

51:2 Look unto Abraham your father, and unto Sarah that nam laci de qua praecisi estis 51:2 adtendite ad Abraham

bare you: for I called him alone, and blessed him, and increased him.

patrem uestrum et ad Sarram quae peperit uos quia unum uo-

51:3 For the LORD shall comfort Zion: he will comfort caui eum et benedixi ei et multiplicaui eum 51:3 consolabi-

all her waste places; and he will make her wilderness like Eden, and her desert like the garden of the LORD; joy and gladness shall be found therein, thanksgiving, and the voice of melody.

tur ergo Dominus et Sion consolabitur omnes ruinas eius et ponet desertum eius quasi delicias et solitudinem eius quasi

Prophetia Isaiae

hortum Domini gaudium et laetitia inuenietur in ea gratiarum actio et uox laudis 51:4 adtendite ad me populus meus 51:4 Hearken unto me, my people; and give ear unto me,

O my nation: for a law shall proceed from me, and I will make my judgment to rest for a light of the people.

et tribus mea me audite quia lex a me exiet et iudicium meum

in lucem populorum requiescet 51:5 prope est iustus meus 51:5 My righteousness is near; my salvation is gone forth, and mine arms shall judge the people; the isles shall wait upon me, and on mine arm shall they trust.

egressus est saluator meus et brachia mea populos iudicabunt me insulae expectabunt et brachium meum sustinebunt 51:6 leuate in caelum oculos uestros et uidete sub terra deor- 51:6 Lift up your eyes to the heavens, and look upon

the earth beneath: for the heavens shall vanish away like smoke, and the earth shall wax old like a garment, and they that dwell therein shall die in like manner: but my salvation shall be for ever, and my righteousness shall not be abolished.

sum quia caeli sicut fumus liquescent et terra sicut uestimentum adteretur et habitatores eius sicut haec interibunt salus autem mea in sempiternum erit et iustitia mea non deficiet 51:7 audite me qui scitis iustum populus lex mea in corde 51:7 Hearken unto me, ye that know righteousness, the

people in whose heart is my law; fear ye not the reproach of men, neither be ye afraid of their revilings.

eorum nolite timere obprobrium hominum et blasphemias eorum ne metuatis 51:8 sicut enim uestimentum sic come- 51:8 For the moth shall eat them up like a garment, and

the worm shall eat them like wool: but my righteousness shall be for ever, and my salvation from generation to gen-

det eos uermis et sicut lanam sic deuorabit eos tinea salus eration. autem mea in sempiternum erit et iustitia mea in generationes generationum 51:9 consurge consurge induere fortitu- 51:9 Awake, awake, put on strength, O arm of the LORD;

awake, as in the ancient days, in the generations of old. Art thou not it that hath cut Rahab, and wounded the dragon?

dinem brachium Domini consurge sicut in diebus antiquis in generationibus saeculorum numquid non tu percussisti superbum uulnerasti draconem 51:10 numquid non tu siccasti 51:10 Art thou not it which hath dried the sea, the waters

of the great deep; that hath made the depths of the sea a way for the ransomed to pass over?

mare aquam abyssi uehementis qui posuisti profundum maris uiam ut transirent liberati 51:11 et nunc qui redeepti sunt 51:11 Therefore the redeemed of the LORD shall return,

and come with singing unto Zion; and everlasting joy shall be upon their head: they shall obtain gladness and joy; and

a Domino reuertentur et uenient in Sion laudantes et laetitia sorrow and mourning shall flee away. sempiterna super capita eorum gaudium et laetitiam tenebunt fugiet dolor et gemitus 51:12 ego ego ipse consolabor 51:12 I, even I, am he that comforteth you: who art thou,

that thou shouldest be afraid of a man that shall die, and of the son of man which shall be made as grass;

uos quis tu ut timeres ab homine mortali et a filio hominis qui quasi faenum ita arescet 51:13 et oblitus es Domini fac- 51:13 And forgettest the LORD thy maker, that hath

stretched forth the heavens, and laid the foundations of the earth; and hast feared continually every day because of the fury of the oppressor, as if he were ready to destroy? and where is the fury of the oppressor?

toris tui qui tetendit caelos et fundauit terram et formidasti iugiter tota die a facie furoris eius qui te tribulabat et parauerat ad perdendum ubi nunc est furor tribulantis 51:14 51:14 The captive exile hasteneth that he may be loosed,

and that he should not die in the pit, nor that his bread should fail.

cito ueniet gradiens ad aperiendum et non interficiet usque ad internicionem nec deficiet panis eius 51:15 ego autem 51:15 But I am the LORD thy God, that divided the sea,

whose waves roared: The LORD of hosts is his name.

sum Dominus Deus tuus qui conturbo mare et intumescunt fluctus eius Dominus exercituum nomen meum 51:16 po- 51:16 And I have put my words in thy mouth, and I have

covered thee in the shadow of mine hand, that I may plant the heavens, and lay the foundations of the earth, and say unto Zion, Thou art my people.

Biblia Sacra Vulgata

sui uerba mea in ore tuo et in umbra manus meae protexi te ut plantes caelos et fundes terram et dicas ad Sion populus

51:17 Awake, awake, stand up, O Jerusalem, which hast meus es tu 51:17 eleuare eleuare consurge Hierusalem quae

drunk at the hand of the LORD the cup of his fury; thou hast drunken the dregs of the cup of trembling, and wrung

them out. bibisti de manu Domini calicem irae eius usque ad fundum 51:18 There is none to guide her among all the sons whom calicis soporis bibisti et epotasti usque ad feces 51:18 non

she hath brought forth; neither is there any that taketh her by the hand of all the sons that she hath brought up.

est qui sustentet eam ex omnibus filiis quos genuit et non est qui adprehendat manum eius ex omnibus filiis quos enutriuit

51:19 These two things are come unto thee; who shall 51:19 duo sunt quae occurrerunt tibi quis contristabitur super

be sorry for thee? desolation, and destruction, and the famine, and the sword: by whom shall I comfort thee?

te uastitas et contritio et fames et gladius quis consolabitur

51:20 Thy sons have fainted, they lie at the head of all the te 51:20 filii tui proiecti sunt dormierunt in capite omnium

streets, as a wild bull in a net: they are full of the fury of the LORD, the rebuke of thy God.

uiarum sicut bestia inlaqueata pleni indignatione Domini in-

51:21 Therefore hear now this, thou afflicted, and drunken, crepatione Dei tui 51:21 idcirco audi hoc paupercula et eb-

but not with wine:

51:22 Thus saith thy Lord the LORD, and thy God that ria non a uino 51:22 haec dicit Dominator tuus Dominus et

pleadeth the cause of his people, Behold, I have taken out of thine hand the cup of trembling, even the dregs of the

cup of my fury; thou shalt no more drink it again: Deus tuus qui pugnauit pro populo suo ecce tuli de manu

tua calicem soporis fundum calicis indignationis meae non

51:23 But I will put it into the hand of them that afflict adicies ut bibas illud ultra 51:23 et ponam illud in manu eo-

thee; which have said to thy soul, Bow down, that we may go over: and thou hast laid thy body as the ground, and as

the street, to them that went over. rum qui te humiliauerunt et dixerunt animae tuae incuruare ut transeamus et posuisti ut terram corpus tuum et quasi uiam transeuntibus

52:1 Awake, awake; put on thy strength, O Zion; put on 52:1 consurge consurge induere fortitudine tua Sion in-

thy beautiful garments, O Jerusalem, the holy city: for henceforth there shall no more come into thee the uncir-

cumcised and the unclean. duere uestimentis gloriae tuae Hierusalem ciuitas sancti quia

non adiciet ultra ut pertranseat per te incircumcisus et in-

52:2 Shake thyself from the dust; arise, and sit down, O mundus 52:2 excutere de puluere consurge sede Hierusa-

Jerusalem: loose thyself from the bands of thy neck, O captive daughter of Zion.

lem solue uincula colli tui captiua filia Sion 52:3 quia haec

52:3 For thus saith the LORD, Ye have sold yourselves for nought; and ye shall be redeemed without money.

dicit Dominus gratis uenundati estis et sine argento rediie-

52:4 For thus saith the Lord GOD, My people went down mini 52:4 quia haec dicit Dominus Deus in Aegyptum de-

aforetime into Egypt to sojourn there; and the Assyrian oppressed them without cause.

scendit populus meus in principio ut colonus esset ibi et

52:5 Now therefore, what have I here, saith the LORD, Assur absque ulla causa calumniatus est eum 52:5 et nunc

that my people is taken away for nought? they that rule over them make them to howl, saith the LORD; and my

name continually every day is blasphemed. quid mihi est hic dicit Dominus quoniam ablatus est popu-

lus meus gratis dominatores eius inique agunt dicit Dominus

52:6 Therefore my people shall know my name: therefore et iugiter tota die nomen meum blasphematur 52:6 propter

they shall know in that day that I am he that doth speak: behold, it is I.

hoc sciet populus meus nomen meum in die illa quia ego ipse

52:7 How beautiful upon the mountains are the feet of qui loquebar ecce adsum 52:7 quam pulchri super montes

him that bringeth good tidings, that publisheth peace; that bringeth good tidings of good, that publisheth salvation; that saith unto Zion, Thy God reigneth!

Prophetia Isaiae

pedes adnuntiantis et praedicantis pacem adnuntiantis bonum praedicantis salutem dicentis Sion regnauit Deus tuus 52:8 uox speculatorum tuorum leuauerunt uocem simul lau- 52:8 Thy watchmen shall lift up the voice; with the voice

together shall they sing: for they shall see eye to eye, when the LORD shall bring again Zion.

dabunt quia oculum ad oculum uidebunt cum conuerterit

Dominus Sion 52:9 gaudete et laudate simul deserta Hie- 52:9 Break forth into joy, sing together, ye waste places of Jerusalem: for the LORD hath comforted his people, he hath redeemed Jerusalem.

rusalem quia consolatus est Dominus populum suum redeeit Hierusalem 52:10 parauit Dominus brachium sanctum 52:10 The LORD hath made bare his holy arm in the eyes

of all the nations; and all the ends of the earth shall see the salvation of our God.

suum in oculis omnium gentium et uidebunt omnes fines terrae salutare Dei nostri 52:11 recedite recedite exite in- 52:11 Depart ye, depart ye, go ye out from thence, touch

no unclean thing; go ye out of the midst of her; be ye clean, that bear the vessels of the LORD.

de pollutum nolite tangere exite de medio eius mundamini qui fertis uasa Domini 52:12 quoniam non in tumultu exibi- 52:12 For ye shall not go out with haste, nor go by flight:

for the LORD will go before you; and the God of Israel will be your rereward.

tis nec in fuga properabitis praecedet enim uos Dominus et congregabit uos Deus Israhel 52:13 ecce intelleget seruus 52:13 Behold, my servant shall deal prudently, he shall be

exalted and extolled, and be very high.

meus exaltabitur et eleuabitur et sublimis erit ualde 52:14 52:14 As many were astonied at thee; his visage was so

marred more than any man, and his form more than the sons of men:

sicut obstipuerunt super te multi sic inglorius erit inter uiros aspectus eius et forma eius inter filios hominum 52:15 52:15 So shall he sprinkle many nations; the kings shall

shut their mouths at him: for that which had not been told them shall they see; and that which they had not heard

iste asperget gentes multas super ipsum continebunt reges os shall they consider. suum quia quibus non est narratum de eo uiderunt et qui non audierunt contemplati sunt

53:1 quis credidit auditui nostro et brachium Domini cui 53:1 Who hath believed our report? and to whom is the

arm of the LORD revealed?

reuelatum est 53:2 et ascendet sicut uirgultum coram eo et 53:2 For he shall grow up before him as a tender plant,

and as a root out of a dry ground: he hath no form nor comeliness; and when we shall see him, there is no beauty

sicut radix de terra sitienti non est species ei neque decor that we should desire him. et uidimus eum et non erat aspectus et desiderauimus eum

53:3 despectum et nouissimum uirorum uirum dolorum et 53:3 He is despised and rejected of men; a man of sorrows, and acquainted with grief: and we hid as it were our faces from him; he was despised, and we esteemed him not.

scientem infirmitatem et quasi absconditus uultus eius et despectus unde nec reputauimus eum 53:4 uere languores no- 53:4 Surely he hath borne our griefs, and carried our sor-

rows: yet we did esteem him stricken, smitten of God, and afflicted.

stros ipse tulit et dolores nostros ipse portauit et nos putauimus eum quasi leprosum et percussum a Deo et humiliatum 53:5 ipse autem uulneratus est propter iniquitates nostras ad- 53:5 But he was wounded for our transgressions, he was

bruised for our iniquities: the chastisement of our peace was upon him; and with his stripes we are healed.

tritus est propter scelera nostra disciplina pacis nostrae super eum et liuore eius sanati sumus 53:6 omnes nos quasi oues 53:6 All we like sheep have gone astray; we have turned

every one to his own way; and the LORD hath laid on him the iniquity of us all.

errauimus unusquisque in uiam suam declinauit et Dominus posuit in eo iniquitatem omnium nostrum 53:7 oblatus est 53:7 He was oppressed, and he was afflicted, yet he

opened not his mouth: he is brought as a lamb to the slaughter, and as a sheep before her shearers is dumb, so he openeth not his mouth.

Biblia Sacra Vulgata

quia ipse uoluit et non aperuit os suum sicut ouis ad occisionem ducetur et quasi agnus coram tondente obmutescet

53:8 He was taken from prison and from judgment: and et non aperiet os suum 53:8 de angustia et de iudicio sub-

who shall declare his generation? for he was cut off out of the land of the living: for the transgression of my people

was he stricken. latus est generationem eius quis enarrabit quia abscisus est

de terra uiuentium propter scelus populi mei percussit eum

53:9 And he made his grave with the wicked, and with the 53:9 et dabit impios pro sepultura et diuitem pro morte sua eo

rich in his death; because he had done no violence, neither was any deceit in his mouth.

quod iniquitatem non fecerit neque dolus fuerit in ore eius

53:10 Yet it pleased the LORD to bruise him; he hath put 53:10 et Dominus uoluit conterere eum in infirmitate si po-

him to grief: when thou shalt make his soul an offering for sin, he shall see his seed, he shall prolong his days, and

the pleasure of the LORD shall prosper in his hand. suerit pro peccato animam suam uidebit semen longeuum et

53:11 He shall see of the travail of his soul, and shall be uoluntas Domini in manu eius dirigetur 53:11 pro eo quod satisfied: by his knowledge shall my righteous servant justify many; for he shall bear their iniquities.

laborauit anima eius uidebit et saturabitur in scientia sua iustificabit ipse iustus seruus meus multos et iniquitates eorum

53:12 Therefore will I divide him a portion with the great, ipse portabit 53:12 ideo dispertiam ei plurimos et fortium

and he shall divide the spoil with the strong; because he hath poured out his soul unto death: and he was numbered with the transgressors; and he bare the sin of many, and made intercession for the transgressors.

diuidet spolia pro eo quod tradidit in morte animam suam et cum sceleratis reputatus est et ipse peccatum multorum tulit et pro transgressoribus rogauit

54:1 Sing, O barren, thou that didst not bear; break forth 54:1 lauda sterilis quae non paris decanta laudem et hinni

into singing, and cry aloud, thou that didst not travail with child: for more are the children of the desolate than the

children of the married wife, saith the LORD. quae non pariebas quoniam multi filii desertae magis quam 54:2 Enlarge the place of thy tent, and let them stretch eius quae habebat uirum dicit Dominus 54:2 dilata locum

forth the curtains of thine habitations: spare not, lengthen thy cords, and strengthen thy stakes;

tentorii tui et pelles tabernaculorum tuorum extende ne par-

54:3 For thou shalt break forth on the right hand and on cas longos fac funiculos tuos et clauos tuos consolida 54:3

the left; and thy seed shall inherit the Gentiles, and make the desolate cities to be inhabited.

ad dexteram enim et ad leuam penetrabis et semen tuum gen-

54:4 Fear not; for thou shalt not be ashamed: neither be tes hereditabit et ciuitates desertas inhabitabit 54:4 noli ti-

thou confounded; for thou shalt not be put to shame: for thou shalt forget the shame of thy youth, and shalt not re-

member the reproach of thy widowhood any more. mere quia non confunderis neque erubescas non enim te pu-

debit quia confusionis adulescentiae tuae obliuisceris et ob-

54:5 For thy Maker is thine husband; the LORD of hosts is probrii uiduitatis tuae non recordaberis amplius 54:5 quia

his name; and thy Redeemer the Holy One of Israel; The God of the whole earth shall he be called.

dominabitur tui qui fecit te Dominus exercituum nomen eius et redeeptor tuus Sanctus Israhel Deus omnis terrae uoca-

54:6 For the LORD hath called thee as a woman forsaken bitur 54:6 quia ut mulierem derelictam et maerentem spi-

and grieved in spirit, and a wife of youth, when thou wast refused, saith thy God.

ritu uocauit te Dominus et uxorem ab adulescentia abiec-

54:7 For a small moment have I forsaken thee; but with tam dixit Deus tuus 54:7 ad punctum in modico dereliqui

great mercies will I gather thee. 54:8 In a little wrath I hid my face from thee for a moment; te et in miserationibus magnis congregabo te 54:8 in mobut with everlasting kindness will I have mercy on thee, saith the LORD thy Redeemer.

mento indignationis abscondi faciem meam parumper a te et

Prophetia Isaiae

in misericordia sempiterna misertus sum tui dixit redeeptor tuus Dominus 54:9 sicut in diebus Noe istud mihi est cui 54:9 For this is as the waters of Noah unto me: for as I

have sworn that the waters of Noah should no more go over the earth; so have I sworn that I would not be wroth

iuraui ne inducerem aquas Noe ultra super terram sic iuraui with thee, nor rebuke thee.

ut non irascar tibi et non increpem te 54:10 montes enim 54:10 For the mountains shall depart, and the hills be removed; but my kindness shall not depart from thee, neither shall the covenant of my peace be removed, saith the

commouebuntur et colles contremescent misericordia autem LORD that hath mercy on thee. mea non recedet et foedus pacis meae non mouebitur dixit miserator tuus Dominus 54:11 paupercula tempestate con- 54:11 O thou afflicted, tossed with tempest, and not com-

forted, behold, I will lay thy stones with fair colours, and lay thy foundations with sapphires.

uulsa absque ulla consolatione ecce ego sternam per ordinem lapides tuos et fundabo te in sapphyris 54:12 et ponam ias- 54:12 And I will make thy windows of agates, and thy

gates of carbuncles, and all thy borders of pleasant stones.

pidem propugnacula tua et portas tuas in lapides sculptos et omnes terminos tuos in lapides desiderabiles 54:13 uniuer- 54:13 And all thy children shall be taught of the LORD;

and great shall be the peace of thy children.

sos filios tuos doctos a Domino et multitudinem pacis filiis tuis 54:14 et in iustitia fundaberis recede procul a calum- 54:14 In righteousness shalt thou be established: thou

shalt be far from oppression; for thou shalt not fear: and from terror; for it shall not come near thee.

nia quia non timebis et a pauore quia non adpropinquabit

tibi 54:15 ecce accola ueniet qui non erat mecum aduena 54:15 Behold, they shall surely gather together, but not by me: whosoever shall gather together against thee shall fall for thy sake.

quondam tuus adiungetur tibi 54:16 ecce ego creaui fabrum

54:16 Behold, I have created the smith that bloweth the coals in the fire, and that bringeth forth an instrument for his work; and I have created the waster to destroy.

sufflantem in igne prunas et proferentem uas in opus suum et

ego creaui interfectorem ad disperdendum 54:17 omne uas 54:17 No weapon that is formed against thee shall prosper; and every tongue that shall rise against thee in judgment thou shalt condemn. This is the heritage of the servants of the LORD, and their righteousness is of me, saith the LORD.

quod fictum est contra te non dirigetur et omnem linguam resistentem tibi in iudicio iudicabis haec hereditas seruorum Domini et iustitia eorum apud me dicit Dominus

55:1 o omnes sitientes uenite ad aquas et qui non habe- 55:1 Ho, every one that thirsteth, come ye to the wa-

ters, and he that hath no money; come ye, buy, and eat; yea, come, buy wine and milk without money and without

tis argentum properate emite et comedite uenite emite abs- price. que argento et absque ulla commutatione uinum et lac 55:2 55:2 Wherefore do ye spend money for that which is not

bread? and your labour for that which satisfieth not? hearken diligently unto me, and eat ye that which is good, and

quare adpenditis argentum non in panibus et laborem ue- let your soul delight itself in fatness. strum non in saturitate audite audientes me et comedite bonum et delectabitur in crassitudine anima uestra 55:3 incli- 55:3 Incline your ear, and come unto me: hear, and your

soul shall live; and I will make an everlasting covenant with you, even the sure mercies of David.

nate aurem uestram et uenite ad me audite et uiuet anima uestra et feriam uobis pactum sempiternum misericordias

Dauid fideles 55:4 ecce testem populis dedi eum ducem ac 55:4 Behold, I have given him for a witness to the people,

a leader and commander to the people.

praeceptorem gentibus 55:5 ecce gentem quam nesciebas 55:5 Behold, thou shalt call a nation that thou knowest not, and nations that knew not thee shall run unto thee because of the LORD thy God, and for the Holy One of Israel; for

uocabis et gentes quae non cognouerunt te ad te current pro- he hath glorified thee. pter Dominum Deum tuum et Sanctum Israhel quia glorifi-

Biblia Sacra Vulgata

55:6 Seek ye the LORD while he may be found, call ye cauit te 55:6 quaerite Dominum dum inueniri potest inuo-

upon him while he is near: 55:7 Let the wicked forsake his way, and the unrighteous cate eum dum prope est 55:7 derelinquat impius uiam suam man his thoughts: and let him return unto the LORD, and he will have mercy upon him; and to our God, for he will

abundantly pardon. et uir iniquus cogitationes suas et reuertatur ad Dominum et

miserebitur eius et ad Deum nostrum quoniam multus est

55:8 For my thoughts are not your thoughts, neither are ad ignoscendum 55:8 non enim cogitationes meae cogita-

your ways my ways, saith the LORD.

tiones uestrae neque uiae uestrae uiae meae dicit Dominus

55:9 For as the heavens are higher than the earth, so are 55:9 quia sicut exaltantur caeli a terra sic exaltatae sunt uiae

my ways higher than your ways, and my thoughts than your thoughts.

meae a uiis uestris et cogitationes meae a cogitationibus ue-

55:10 For as the rain cometh down, and the snow from stris 55:10 et quomodo descendit imber et nix de caelo et

heaven, and returneth not thither, but watereth the earth, and maketh it bring forth and bud, that it may give seed to

the sower, and bread to the eater: illuc ultra non reuertitur sed inebriat terram et infundit eam

et germinare eam facit et dat semen serenti et panem come-

55:11 So shall my word be that goeth forth out of my denti 55:11 sic erit uerbum meum quod egredietur de ore

mouth: it shall not return unto me void, but it shall accomplish that which I please, and it shall prosper in the

thing whereto I sent it. meo non reuertetur ad me uacuum sed faciet quaecumque 55:12 For ye shall go out with joy, and be led forth with uolui et prosperabitur in his ad quae misi illud 55:12 quia

peace: the mountains and the hills shall break forth before you into singing, and all the trees of the field shall clap

their hands. in laetitia egrediemini et in pace deducemini montes et colles

cantabunt coram uobis laudem et omnia ligna regionis plau-

55:13 Instead of the thorn shall come up the fir tree, and dent manu 55:13 pro saliunca ascendet abies et pro urtica

instead of the brier shall come up the myrtle tree: and it shall be to the LORD for a name, for an everlasting sign

that shall not be cut off. crescet myrtus et erit Dominus nominatus in signum aeternum quod non auferetur

56:1 Thus saith the LORD, Keep ye judgment, and do jus- 56:1 haec dicit Dominus custodite iudicium et facite ius-

tice: for my salvation is near to come, and my righteousness to be revealed.

titiam quia iuxta est salus mea ut ueniat et iustitia mea ut

56:2 Blessed is the man that doeth this, and the son of reueletur 56:2 beatus uir qui facit hoc et filius hominis qui

man that layeth hold on it; that keepeth the sabbath from polluting it, and keepeth his hand from doing any evil.

adprehendit istud custodiens sabbatum ne polluat illud cus-

56:3 Neither let the son of the stranger, that hath joined todiens manus suas ne faciat omne malum 56:3 et non dicat himself to the LORD, speak, saying, The LORD hath utterly separated me from his people: neither let the eunuch

say, Behold, I am a dry tree. filius aduenae qui adheret Domino dicens separatione diui-

det me Dominus a populo suo et non dicat eunuchus ecce

56:4 For thus saith the LORD unto the eunuchs that keep ego lignum aridum 56:4 quia haec dicit Dominus eunuchis

my sabbaths, and choose the things that please me, and take hold of my covenant;

qui custodierint sabbata mea et elegerint quae uolui et tenue-

56:5 Even unto them will I give in mine house and within rint foedus meum 56:5 dabo eis in domo mea et in muris my walls a place and a name better than of sons and of daughters: I will give them an everlasting name, that shall

not be cut off. meis locum et nomen melius a filiis et filiabus nomen sem-

56:6 Also the sons of the stranger, that join themselves piternum dabo eis quod non peribit 56:6 et filios aduenae to the LORD, to serve him, and to love the name of the LORD, to be his servants, every one that keepeth the sab-

bath from polluting it, and taketh hold of my covenant; qui adherent Domino ut colant eum et diligant nomen eius ut sint ei in seruos omnem custodientem sabbatum ne pol-

Prophetia Isaiae

luat illud et tenentem foedus meum 56:7 adducam eos in 56:7 Even them will I bring to my holy mountain, and

make them joyful in my house of prayer: their burnt offerings and their sacrifices shall be accepted upon mine altar; for mine house shall be called an house of prayer for all people.

montem sanctum meum et laetificabo eos in domo orationis meae holocausta eorum et uictimae eorum placebunt mihi super altari meo quia domus mea domus orationis uocabitur cunctis populis 56:8 ait Dominus Deus qui congregat dis- 56:8 The Lord GOD, which gathereth the outcasts of Israel

saith, Yet will I gather others to him, beside those that are gathered unto him.

persos Israhel adhuc congregabo ad eum congregatos eius 56:9 omnes bestiae agri uenite ad deuorandum uniuersae be- 56:9 All ye beasts of the field, come to devour, yea, all ye

beasts in the forest.

stiae saltus 56:10 speculatores eius caeci omnes nescierunt 56:10 His watchmen are blind: they are all ignorant, they

are all dumb dogs, they cannot bark; sleeping, lying down, loving to slumber.

uniuersi canes muti non ualentes latrare uidentes uana dormientes et amantes somnia 56:11 et canes inpudentissimi 56:11 Yea, they are greedy dogs which can never have

enough, and they are shepherds that cannot understand: they all look to their own way, every one for his gain, from

nescierunt saturitatem ipsi pastores ignorauerunt intellegen- his quarter. tiam omnes in uiam suam declinauerunt unusquisque ad aua-

ritiam suam a summo usque ad nouissimum 56:12 uenite 56:12 Come ye, say they, I will fetch wine, and we will fill ourselves with strong drink; and to morrow shall be as this day, and much more abundant.

sumamus uinum et impleamur ebrietate et erit sicut hodie sic et cras et multo amplius

57:1 iustus perit et nemo est qui recogitet in corde suo et 57:1 The righteous perisheth, and no man layeth it to heart: and merciful men are taken away, none considering that the righteous is taken away from the evil to come.

uiri misericordiae colliguntur quia non est qui intellegat a facie enim malitiae collectus est iustus 57:2 ueniat pax re- 57:2 He shall enter into peace: they shall rest in their beds,

each one walking in his uprightness.

quiescat in cubili suo qui ambulauit in directione sua 57:3 57:3 But draw near hither, ye sons of the sorceress, the

seed of the adulterer and the whore.

uos autem accedite huc filii auguratricis semen adulteri et fornicariae 57:4 super quem lusistis super quem dilatastis 57:4 Against whom do ye sport yourselves? against whom

make ye a wide mouth, and draw out the tongue? are ye not children of transgression, a seed of falsehood.

os et eiecistis linguam numquid non uos filii scelesti semen mendax 57:5 qui consolamini in diis subter omne li- 57:5 Enflaming yourselves with idols under every green

tree, slaying the children in the valleys under the clifts of the rocks?

gnum frondosum immolantes paruulos in torrentibus subter inminentes petras 57:6 in partibus torrentis pars tua haec 57:6 Among the smooth stones of the stream is thy por-

tion; they, they are thy lot: even to them hast thou poured a drink offering, thou hast offered a meat offering. Should

est sors tua et ipsis effudisti libamen obtulisti sacrificium I receive comfort in these?

numquid super his non indignabor 57:7 super montem ex- 57:7 Upon a lofty and high mountain hast thou set thy bed:

even thither wentest thou up to offer sacrifice.

celsum et sublimem posuisti cubile tuum et illuc ascendisti ut immolares hostias 57:8 et post ostium et retro postem 57:8 Behind the doors also and the posts hast thou set up

thy remembrance: for thou hast discovered thyself to another than me, and art gone up; thou hast enlarged thy bed, and made thee a covenant with them; thou lovedst their bed where thou sawest it.

posuisti memoriale tuum quia iuxta me discoperuisti et suscepisti adulterum dilatasti cubile tuum et pepigisti cum eis dilexisti stratum eorum manu aperta 57:9 et ornasti te regi 57:9 And thou wentest to the king with ointment, and didst

increase thy perfumes, and didst send thy messengers far off, and didst debase thyself even unto hell.

unguento et multiplicasti pigmenta tua misisti legatos tuos

Biblia Sacra Vulgata

57:10 Thou art wearied in the greatness of thy way; yet procul et humiliata es usque ad inferos 57:10 in multitu-

saidst thou not, There is no hope: thou hast found the life of thine hand; therefore thou wast not grieved.

dine uiae tuae laborasti non dixisti quiescam uitam manus

57:11 And of whom hast thou been afraid or feared, that tuae inuenisti propterea non rogasti 57:11 pro quo sollicita

thou hast lied, and hast not remembered me, nor laid it to thy heart? have not I held my peace even of old, and thou

fearest me not? timuisti quia mentita es et mei non es recordata neque cogi-

tasti in corde tuo quia ego tacens et quasi non uidens et mei

57:12 I will declare thy righteousness, and thy works; for oblita es 57:12 ego adnuntiabo iustitiam tuam et opera tua they shall not profit thee.

57:13 When thou criest, let thy companies deliver thee; non prodeeunt tibi 57:13 cum clamaueris liberent te con-

but the wind shall carry them all away; vanity shall take them: but he that putteth his trust in me shall possess the

land, and shall inherit my holy mountain; gregati tui et omnes eos auferet uentus tollet aura qui autem

fiduciam habet mei hereditabit terram et possidebit mon-

57:14 And shall say, Cast ye up, cast ye up, prepare the tem sanctum meum 57:14 et dicam uiam facite praebete way, take up the stumblingblock out of the way of my people.

iter declinate de semita auferte offendicula de uia populi mei

57:15 For thus saith the high and lofty One that inhabiteth 57:15 quia haec dicit Excelsus et Sublimis habitans aeternita-

eternity, whose name is Holy; I dwell in the high and holy place, with him also that is of a contrite and humble spirit, to revive the spirit of the humble, and to revive the heart of the contrite ones.

tem et sanctum nomen eius in excelso et in sancto habitans et cum contrito et humili spiritu ut uiuificet spiritum humi-

57:16 For I will not contend for ever, neither will I be al- lium et uiuificet cor contritorum 57:16 non enim in sempiways wroth: for the spirit should fail before me, and the souls which I have made.

ternum litigabo neque usque ad finem irascar quia spiritus a

57:17 For the iniquity of his covetousness was I wroth, facie mea egredietur et flatus ego faciam 57:17 propter in-

and smote him: I hid me, and was wroth, and he went on frowardly in the way of his heart.

iquitatem auaritiae eius iratus sum et percussi eum abscondi

57:18 I have seen his ways, and will heal him: I will lead et indignatus sum et abiit uagus in uia cordis sui 57:18 uias

him also, and restore comforts unto him and to his mourners.

eius uidi et dimisi eum et reduxi eum et reddidi consolatio-

57:19 I create the fruit of the lips; Peace, peace to him that nes ipsi et lugentibus eius 57:19 creaui fructum labiorum is far off, and to him that is near, saith the LORD; and I will heal him.

pacem pacem ei qui longe est et qui prope dixit Dominus

57:20 But the wicked are like the troubled sea, when it et sanaui eum 57:20 impii autem quasi mare feruens quod

cannot rest, whose waters cast up mire and dirt.

quiescere non potest et redundant fluctus eius in conculca-

57:21 There is no peace, saith my God, to the wicked. tionem et lutum 57:21 non est pax dixit Deus meus impiis

58:1 Cry aloud, spare not, lift up thy voice like a trumpet, 58:1 clama ne cesses quasi tuba exalta uocem tuam et adand shew my people their transgression, and the house of Jacob their sins.

nuntia populo meo scelera eorum et domui Iacob peccata

58:2 Yet they seek me daily, and delight to know my ways, eorum 58:2 me etenim de die in diem quaerunt et scire as a nation that did righteousness, and forsook not the ordinance of their God: they ask of me the ordinances of

justice; they take delight in approaching to God. uias meas uolunt quasi gens quae iustitiam fecerit et quae

iudicium Dei sui non reliquerit rogant me iudicia iustitiae

58:3 Wherefore have we fasted, say they, and thou seest adpropinquare Deo uolunt 58:3 quare ieiunauimus et non

not? wherefore have we afflicted our soul, and thou takest no knowledge? Behold, in the day of your fast ye find

pleasure, and exact all your labours. aspexisti humiliauimus animam nostram et nescisti ecce in die ieiunii uestri inuenitur uoluntas et omnes debitores ue-

Prophetia Isaiae

stros repetitis 58:4 ecce ad lites et contentiones ieiunatis et 58:4 Behold, ye fast for strife and debate, and to smite

with the fist of wickedness: ye shall not fast as ye do this day, to make your voice to be heard on high.

percutitis pugno impie nolite ieiunare sicut usque ad hanc diem ut audiatur in excelso clamor uester 58:5 numquid 58:5 Is it such a fast that I have chosen? a day for a man

to afflict his soul? is it to bow down his head as a bulrush, and to spread sackcloth and ashes under him? wilt thou

tale est ieiunium quod elegi per diem adfligere hominem call this a fast, and an acceptable day to the LORD? animam suam numquid contorquere quasi circulum caput suum et saccum et cinerem sternere numquid istud uocabis ieiunium et diem acceptabilem Domino 58:6 nonne hoc est 58:6 Is not this the fast that I have chosen? to loose the

bands of wickedness, to undo the heavy burdens, and to let the oppressed go free, and that ye break every yoke?

magis ieiunium quod elegi dissolue conligationes impietatis solue fasciculos deprimentes dimitte eos qui confracti sunt liberos et omne onus disrumpe 58:7 frange esurienti panem 58:7 Is it not to deal thy bread to the hungry, and that thou

bring the poor that are cast out to thy house? when thou seest the naked, that thou cover him; and that thou hide

tuum et egenos uagosque induc in domum tuam cum uideris not thyself from thine own flesh?

nudum operi eum et carnem tuam ne despexeris 58:8 tunc 58:8 Then shall thy light break forth as the morning, and thine health shall spring forth speedily: and thy righteousness shall go before thee; the glory of the LORD shall be

erumpet quasi mane lumen tuum et sanitas tua citius orietur thy rereward. et anteibit faciem tuam iustitia tua et gloria Domini colliget te 58:9 tunc inuocabis et Dominus exaudiet clamabis et 58:9 Then shalt thou call, and the LORD shall answer;

thou shalt cry, and he shall say, Here I am. If thou take away from the midst of thee the yoke, the putting forth of

dicet ecce adsum si abstuleris de medio tui catenam et desie- the finger, and speaking vanity; ris digitum extendere et loqui quod non prodeet 58:10 cum 58:10 And if thou draw out thy soul to the hungry, and sat-

isfy the afflicted soul; then shall thy light rise in obscurity, and thy darkness be as the noon day:

effuderis esurienti animam tuam et animam adflictam repleueris orietur in tenebris lux tua et tenebrae tuae erunt sicut meridies 58:11 et requiem tibi dabit Dominus semper et im- 58:11 And the LORD shall guide thee continually, and sat-

isfy thy soul in drought, and make fat thy bones: and thou shalt be like a watered garden, and like a spring of water,

plebit splendoribus animam tuam et ossa tua liberabit et eris whose waters fail not. quasi hortus inriguus et sicut fons aquarum cuius non deficient aquae 58:12 et aedificabuntur in te deserta saeculorum 58:12 And they that shall be of thee shall build the old

waste places: thou shalt raise up the foundations of many generations; and thou shalt be called, The repairer of the

fundamenta generationis et generationis suscitabis et uoca- breach, The restorer of paths to dwell in. beris aedificator sepium auertens semitas in quietem 58:13 58:13 If thou turn away thy foot from the sabbath, from

doing thy pleasure on my holy day; and call the sabbath a delight, the holy of the LORD, honourable; and shalt honour him, not doing thine own ways, nor finding thine own pleasure, nor speaking thine own words:

si auerteris a sabbato pedem tuum facere uoluntatem tuam in die sancto meo et uocaueris sabbatum delicatum et sanctum Domini gloriosum et glorificaueris eum dum non facis uias tuas et non inuenitur uoluntas tua ut loquaris sermonem 58:14 tunc delectaberis super Domino et sustollam te super 58:14 Then shalt thou delight thyself in the LORD; and I

will cause thee to ride upon the high places of the earth, and feed thee with the heritage of Jacob thy father: for the

altitudines terrae et cibabo te hereditate Iacob patris tui os mouth of the LORD hath spoken it. enim Domini locutum est

59:1 ecce non est adbreuiata manus Domini ut saluare ne- 59:1 Behold, the LORD’s hand is not shortened, that it

cannot save; neither his ear heavy, that it cannot hear:

Biblia Sacra Vulgata

59:2 But your iniquities have separated between you and queat neque adgrauata est auris eius ut non exaudiat 59:2

your God, and your sins have hid his face from you, that he will not hear.

sed iniquitates uestrae diuiserunt inter uos et Deum uestrum et peccata uestra absconderunt faciem eius a uobis ne ex-

59:3 For your hands are defiled with blood, and your fin- audiret 59:3 manus enim uestrae pollutae sunt sanguine et

gers with iniquity; your lips have spoken lies, your tongue hath muttered perverseness.

digiti uestri iniquitate labia uestra locuta sunt mendacium et

59:4 None calleth for justice, nor any pleadeth for truth: lingua uestra iniquitatem fatur 59:4 non est qui inuocet ius-

they trust in vanity, and speak lies; they conceive mischief, and bring forth iniquity.

titiam neque est qui iudicet uere sed confidunt in nihili et loquuntur uanitates conceperunt laborem et pepererunt iniqui-

59:5 They hatch cockatrice’ eggs, and weave the spider’s tatem 59:5 oua aspidum ruperunt et telas araneae texuerunt

web: he that eateth of their eggs dieth, and that which is crushed breaketh out into a viper.

qui comederit de ouis eorum morietur et quod confotum est

59:6 Their webs shall not become garments, neither shall erumpet in regulum 59:6 telae eorum non erunt in uestithey cover themselves with their works: their works are works of iniquity, and the act of violence is in their hands.

mentum neque operientur operibus suis opera eorum opera

59:7 Their feet run to evil, and they make haste to shed inutilia et opus iniquitatis in manibus eorum 59:7 pedes eo-

innocent blood: their thoughts are thoughts of iniquity; wasting and destruction are in their paths.

rum ad malum currunt et festinant ut effundant sanguinem innocentem cogitationes eorum cogitationes inutiles uastitas

59:8 The way of peace they know not; and there is no et contritio in uiis eorum 59:8 uiam pacis nescierunt et non

judgment in their goings: they have made them crooked paths: whosoever goeth therein shall not know peace.

est iudicium in gressibus eorum semitae eorum incuruatae

59:9 Therefore is judgment far from us, neither doth jus- sunt eis omnis qui calcat in ea ignorat pacem 59:9 propter tice overtake us: we wait for light, but behold obscurity; for brightness, but we walk in darkness.

hoc elongatum est iudicium a nobis et non adprehendet nos iustitia expectauimus lucem et ecce tenebrae splendorem et

59:10 We grope for the wall like the blind, and we grope in tenebris ambulauimus 59:10 palpauimus sicut caeci pa-

as if we had no eyes: we stumble at noon day as in the night; we are in desolate places as dead men.

rietem et quasi absque oculis adtrectauimus inpegimus me-

59:11 We roar all like bears, and mourn sore like doves: ridie quasi in tenebris in caligosis quasi mortui 59:11 ru-

we look for judgment, but there is none; for salvation, but it is far off from us.

giemus quasi ursi omnes et quasi columbae meditantes gememus expectauimus iudicium et non est salutem et elon-

59:12 For our transgressions are multiplied before thee, gata est a nobis 59:12 multiplicatae sunt enim iniquitates

and our sins testify against us: for our transgressions are with us; and as for our iniquities, we know them;

nostrae coram te et peccata nostra responderunt nobis quia scelera nostra nobiscum et iniquitates nostras cognouimus

59:13 In transgressing and lying against the LORD, and 59:13 peccare et mentiri contra Dominum et auersi sumus ne

departing away from our God, speaking oppression and revolt, conceiving and uttering from the heart words of

falsehood. iremus post tergum Dei nostri ut loqueremur calumniam et

transgressionem concepimus et locuti sumus de corde uerba

59:14 And judgment is turned away backward, and justice mendacii 59:14 et conuersum est retrorsum iudicium et ius-

standeth afar off: for truth is fallen in the street, and equity cannot enter.

titia longe stetit quia corruit in platea ueritas et aequitas non

Prophetia Isaiae

potuit ingredi 59:15 et facta est ueritas in obliuione et qui 59:15 Yea, truth faileth; and he that departeth from evil

maketh himself a prey: and the LORD saw it, and it displeased him that there was no judgment.

recessit a malo praedae patuit et uidit Dominus et malum apparuit in oculis eius quia non est iudicium 59:16 et uidit 59:16 And he saw that there was no man, and wondered

that there was no intercessor: therefore his arm brought salvation unto him; and his righteousness, it sustained

quia non est uir et aporiatus est quia non est qui occurrat et him. saluauit sibi brachium suum et iustitia eius ipsa confirmauit eum 59:17 indutus est iustitia ut lorica et galea salutis 59:17 For he put on righteousness as a breastplate, and an

helmet of salvation upon his head; and he put on the garments of vengeance for clothing, and was clad with zeal

in capite eius indutus est uestimentis ultionis et opertus est as a cloak. quasi pallio zeli 59:18 sicut ad uindictam quasi ad retribu- 59:18 According to their deeds, accordingly he will repay,

fury to his adversaries, recompence to his enemies; to the islands he will repay recompence.

tionem indignationis hostibus suis et uicissitudinem inimicis suis insulis uicem reddet 59:19 et timebunt qui ab occidente 59:19 So shall they fear the name of the LORD from the

west, and his glory from the rising of the sun. When the enemy shall come in like a flood, the Spirit of the LORD

nomen Domini et qui ab ortu solis gloriam eius cum uenerit shall lift up a standard against him. quasi fluuius uiolentus quem spiritus Domini cogit 59:20 et 59:20 And the Redeemer shall come to Zion, and unto

them that turn from transgression in Jacob, saith the LORD.

uenerit Sion redeeptor et eis qui redeent ab iniquitate in Ia-

cob dicit Dominus 59:21 hoc foedus meum cum eis dicit 59:21 As for me, this is my covenant with them, saith the LORD; My spirit that is upon thee, and my words which I have put in thy mouth, shall not depart out of thy mouth, nor out of the mouth of thy seed, nor out of the mouth of thy seed’s seed, saith the LORD, from henceforth and for ever.

Dominus spiritus meus qui est in te et uerba mea quae posui in ore tuo non recedent de ore tuo et de ore seminis tui et de ore seminis seminis tui dixit Dominus amodo et usque in sempiternum

60:1 surge inluminare quia uenit lumen tuum et gloria Do- 60:1 Arise, shine; for thy light is come, and the glory of

the LORD is risen upon thee.

mini super te orta est 60:2 quia ecce tenebrae operient ter- 60:2 For, behold, the darkness shall cover the earth, and

gross darkness the people: but the LORD shall arise upon thee, and his glory shall be seen upon thee.

ram et caligo populos super te autem orietur Dominus et gloria eius in te uidebitur 60:3 et ambulabunt gentes in lumine 60:3 And the Gentiles shall come to thy light, and kings to

the brightness of thy rising.

tuo et reges in splendore ortus tui 60:4 leua in circuitu ocu- 60:4 Lift up thine eyes round about, and see: all they

gather themselves together, they come to thee: thy sons shall come from far, and thy daughters shall be nursed at

los tuos et uide omnes isti congregati sunt uenerunt tibi filii thy side. tui de longe uenient et filiae tuae in latere sugent 60:5 tunc 60:5 Then thou shalt see, and flow together, and thine heart

shall fear, and be enlarged; because the abundance of the sea shall be converted unto thee, the forces of the Gentiles

uidebis et afflues et mirabitur et dilatabitur cor tuum quando shall come unto thee. conuersa fuerit ad te multitudo maris fortitudo gentium uenerit tibi 60:6 inundatio camelorum operiet te dromedariae 60:6 The multitude of camels shall cover thee, the

dromedaries of Midian and Ephah; all they from Sheba shall come: they shall bring gold and incense; and they

Madian et Efa omnes de Saba uenient aurum et tus deferen- shall shew forth the praises of the LORD. tes et laudem Domino adnuntiantes 60:7 omne pecus Cedar 60:7 All the flocks of Kedar shall be gathered together

unto thee, the rams of Nebaioth shall minister unto thee: they shall come up with acceptance on mine altar, and I

congregabitur tibi arietes Nabaioth ministrabunt tibi offeren- will glorify the house of my glory. tur super placabili altari meo et domum maiestatis meae glorificabo 60:8 qui sunt isti qui ut nubes uolant et quasi co- 60:8 Who are these that fly as a cloud, and as the doves to

their windows?

Biblia Sacra Vulgata

60:9 Surely the isles shall wait for me, and the ships of lumbae ad fenestras suas 60:9 me enim insulae expectant

Tarshish first, to bring thy sons from far, their silver and their gold with them, unto the name of the LORD thy God, and to the Holy One of Israel, because he hath glorified thee.

et naues maris in principio ut adducam filios tuos de longe argentum eorum et aurum eorum cum eis nomini Domini

60:10 And the sons of strangers shall build up thy walls, Dei tui et Sancto Israhel quia glorificauit te 60:10 et aedi-

and their kings shall minister unto thee: for in my wrath I smote thee, but in my favour have I had mercy on thee.

ficabunt filii peregrinorum muros tuos et reges eorum ministrabunt tibi in indignatione enim mea percussi te et in re-

60:11 Therefore thy gates shall be open continually; they conciliatione mea misertus sum tui 60:11 et aperientur por-

shall not be shut day nor night; that men may bring unto thee the forces of the Gentiles, and that their kings may be

brought. tae tuae iugiter die et nocte non claudentur ut adferatur ad 60:12 For the nation and kingdom that will not serve thee te fortitudo gentium et reges earum adducantur 60:12 gens

shall perish; yea, those nations shall be utterly wasted.

enim et regnum quod non seruierit tibi peribit et gentes so-

60:13 The glory of Lebanon shall come unto thee, the fir litudine uastabuntur 60:13 gloria Libani ad te ueniet abies

tree, the pine tree, and the box together, to beautify the place of my sanctuary; and I will make the place of my

feet glorious. et buxus et pinus simul ad ornandum locum sanctificationis

60:14 The sons also of them that afflicted thee shall come meae et locum pedum meorum glorificabo 60:14 et uenient bending unto thee; and all they that despised thee shall bow themselves down at the soles of thy feet; and they shall call thee; The city of the LORD, The Zion of the Holy One of Israel.

ad te curui filii eorum qui humiliauerunt te et adorabunt uestigia pedum tuorum omnes qui detrahebant tibi et uocabunt

60:15 Whereas thou has been forsaken and hated, so that te ciuitatem Domini Sion Sancti Israhel 60:15 pro eo quod

no man went through thee, I will make thee an eternal excellency, a joy of many generations.

fuisti derelicta et odio habita et non erat qui per te transiret ponam te in superbiam saeculorum gaudium in generatio-

60:16 Thou shalt also suck the milk of the Gentiles, and nem et generationem 60:16 et suges lac gentium et mamilla

shalt suck the breast of kings: and thou shalt know that I the LORD am thy Saviour and thy Redeemer, the mighty

One of Jacob. regum lactaberis et scies quia ego Dominus saluans te et red-

60:17 For brass I will bring gold, and for iron I will bring eeptor tuus Fortis Iacob 60:17 pro aere adferam aurum et silver, and for wood brass, and for stones iron: I will also make thy officers peace, and thine exactors righteousness.

pro ferro adferam argentum et pro lignis aes et pro lapidibus ferrum et ponam uisitationem tuam pacem et praepositos

60:18 Violence shall no more be heard in thy land, wasting tuos iustitiam 60:18 non audietur ultra iniquitas in terra tua

nor destruction within thy borders; but thou shalt call thy walls Salvation, and thy gates Praise.

uastitas et contritio in terminis tuis et occupabit salus muros

60:19 The sun shall be no more thy light by day; neither tuos et portas tuas laudatio 60:19 non erit tibi amplius sol

for brightness shall the moon give light unto thee: but the LORD shall be unto thee an everlasting light, and thy God

thy glory. ad lucendum per diem nec splendor lunae inluminabit te sed

erit tibi Dominus in lucem sempiternam et Deus tuus in glo-

60:20 Thy sun shall no more go down; neither shall thy riam tuam 60:20 non occidet ultra sol tuus et luna tua non

moon withdraw itself: for the LORD shall be thine everlasting light, and the days of thy mourning shall be ended.

minuetur quia Dominus erit in lucem sempiternam et con-

60:21 Thy people also shall be all righteous: they shall plebuntur dies luctus tui 60:21 populus autem tuus omnes

inherit the land for ever, the branch of my planting, the work of my hands, that I may be glorified.

iusti in perpetuum hereditabunt terram germen plantationis

60:22 A little one shall become a thousand, and a small meae opus manus meae ad glorificandum 60:22 minimus one a strong nation: I the LORD will hasten it in his time.

Prophetia Isaiae

erit in mille et paruulus in gentem fortissimam ego Dominus in tempore eius subito faciam istud

61:1 spiritus Domini super me eo quod unxerit Dominus 61:1 The Spirit of the Lord GOD is upon me; because the

LORD hath anointed me to preach good tidings unto the meek; he hath sent me to bind up the brokenhearted, to proclaim liberty to the captives, and the opening of the prison to them that are bound;

me ad adnuntiandum mansuetis misit me ut mederer contritis corde et praedicarem captiuis indulgentiam et clausis apertionem 61:2 ut praedicarem annum placabilem Do- 61:2 To proclaim the acceptable year of the LORD, and the

day of vengeance of our God; to comfort all that mourn;

mini et diem ultionis Deo nostro ut consolarer omnes lugentes 61:3 ut ponerem lugentibus Sion et darem eis coronam 61:3 To appoint unto them that mourn in Zion, to give unto

them beauty for ashes, the oil of joy for mourning, the garment of praise for the spirit of heaviness; that they might be called trees of righteousness, the planting of the LORD, that he might be glorified.

pro cinere oleum gaudii pro luctu pallium laudis pro spiritu maeroris et uocabuntur in ea fortes iustitiae plantatio Do-

mini ad glorificandum 61:4 et aedificabunt deserta a sae- 61:4 And they shall build the old wastes, they shall raise up the former desolations, and they shall repair the waste cities, the desolations of many generations.

culo et ruinas antiquas erigent et instaurabunt ciuitates desertas dissipatas in generationem et generationem 61:5 et 61:5 And strangers shall stand and feed your flocks, and

the sons of the alien shall be your plowmen and your vinedressers.

stabunt alieni et pascent pecora uestra et filii peregrinorum agricolae et uinitores uestri erunt 61:6 uos autem sacerdo- 61:6 But ye shall be named the Priests of the LORD: men

shall call you the Ministers of our God: ye shall eat the riches of the Gentiles, and in their glory shall ye boast

tes Domini uocabimini ministri Dei nostri dicetur uobis for- yourselves. titudinem gentium comedetis et in gloria earum superbietis 61:7 pro confusione uestra duplici et rubore laudabunt 61:7 For your shame ye shall have double; and for confu-

sion they shall rejoice in their portion: therefore in their land they shall possess the double: everlasting joy shall be

partem eorum propter hoc in terra sua duplicia possidebunt unto them. laetitia sempiterna erit eis 61:8 quia ego Dominus diligens 61:8 For I the LORD love judgment, I hate robbery for

burnt offering; and I will direct their work in truth, and I will make an everlasting covenant with them.

iudicium odio habens rapinam in holocausto et dabo opus eorum in ueritate et foedus perpetuum feriam eis 61:9 et 61:9 And their seed shall be known among the Gentiles,

and their offspring among the people: all that see them shall acknowledge them, that they are the seed which the

scietur in gentibus semen eorum et germen eorum in medio LORD hath blessed. populorum omnes qui uiderint eos cognoscent eos quia isti sunt semen cui benedixit Dominus 61:10 gaudens gaudebo 61:10 I will greatly rejoice in the LORD, my soul shall

be joyful in my God; for he hath clothed me with the garments of salvation, he hath covered me with the robe of righteousness, as a bridegroom decketh himself with ornaments, and as a bride adorneth herself with her jewels.

in Domino et exultabit anima mea in Deo meo quia induit me uestimentis salutis et indumento iustitiae circumdedit me quasi sponsum decoratum corona et quasi sponsam ornatam monilibus suis 61:11 sicut enim terra profert germen suum 61:11 For as the earth bringeth forth her bud, and as the

garden causeth the things that are sown in it to spring forth; so the Lord GOD will cause righteousness and praise to

et sicut hortus semen suum germinat sic Dominus Deus ger- spring forth before all the nations. minabit iustitiam et laudem coram uniuersis gentibus

62:1 propter Sion non tacebo et propter Hierusalem non 62:1 For Zion’s sake will I not hold my peace, and for

Jerusalem’s sake I will not rest, until the righteousness thereof go forth as brightness, and the salvation thereof

quiescam donec egrediatur ut splendor iustus eius et saluator as a lamp that burneth.

Biblia Sacra Vulgata

62:2 And the Gentiles shall see thy righteousness, and all eius ut lampas accendatur 62:2 et uidebunt gentes iustum kings thy glory: and thou shalt be called by a new name, which the mouth of the LORD shall name.

tuum et cuncti reges inclitum tuum et uocabitur tibi nomen

62:3 Thou shalt also be a crown of glory in the hand of the nouum quod os Domini nominabit 62:3 et eris corona glo-LORD, and a royal diadem in the hand of thy God. 62:4 Thou shalt no more be termed Forsaken; neither shall riae in manu Domini et diadema regni in manu Dei tui 62:4 thy land any more be termed Desolate: but thou shalt be called Hephzibah, and thy land Beulah: for the LORD de-

lighteth in thee, and thy land shall be married. non uocaberis ultra Derelicta et terra tua non uocabitur amplius Desolata sed uocaberis Voluntas mea in ea et terra tua Inhabitata quia conplacuit Domino in te et terra tua inhabita-

62:5 For as a young man marrieth a virgin, so shall thy bitur 62:5 habitabit enim iuuenis cum uirgine et habitabunt

sons marry thee: and as the bridegroom rejoiceth over the bride, so shall thy God rejoice over thee.

in te filii tui et gaudebit sponsus super sponsam gaudebit

62:6 I have set watchmen upon thy walls, O Jerusalem, super te Deus tuus 62:6 super muros tuos Hierusalem con-

which shall never hold their peace day nor night: ye that make mention of the LORD, keep not silence,

stitui custodes tota die et tota nocte perpetuo non tacebunt

62:7 And give him no rest, till he establish, and till he qui reminiscimini Domini ne taceatis 62:7 et ne detis silen-

make Jerusalem a praise in the earth.

tium ei donec stabiliat et donec ponat Hierusalem laudem

62:8 The LORD hath sworn by his right hand, and by the in terra 62:8 iurauit Dominus in dextera sua et in brachio

arm of his strength, Surely I will no more give thy corn to be meat for thine enemies; and the sons of the stranger

shall not drink thy wine, for the which thou hast laboured: fortitudinis suae si dedero triticum tuum ultra cibum inimi-

cis tuis et si biberint filii alieni uinum tuum in quo laborasti

62:9 But they that have gathered it shall eat it, and praise 62:9 quia qui congregabunt illud comedent et laudabunt Do-

the LORD; and they that have brought it together shall drink it in the courts of my holiness.

minum et qui conportant illud bibent in atriis sanctis meis

62:10 Go through, go through the gates; prepare ye the 62:10 transite transite per portas praeparate uiam populo pla-

way of the people; cast up, cast up the highway; gather out the stones; lift up a standard for the people.

num facite iter et eligite lapides eleuate signum ad populos

62:11 Behold, the LORD hath proclaimed unto the end 62:11 ecce Dominus auditum fecit in extremis terrae dicite fi-

of the world, Say ye to the daughter of Zion, Behold, thy salvation cometh; behold, his reward is with him, and his

work before him. liae Sion ecce saluator tuus uenit ecce merces eius cum eo et 62:12 And they shall call them, The holy people, The re- opus eius coram illo 62:12 et uocabunt eos Populus sanctus

deemed of the LORD: and thou shalt be called, Sought out, A city not forsaken.

Redeepti a Domino tu autem uocaberis Quaesita ciuitas et non Derelicta

63:1 Who is this that cometh from Edom, with dyed gar- 63:1 quis est iste qui uenit de Edom tinctis uestibus de

ments from Bozrah? this that is glorious in his apparel, travelling in the greatness of his strength? I that speak in

righteousness, mighty to save. Bosra iste formonsus in stola sua gradiens in multitudine for-

titudinis suae ego qui loquor iustitiam et propugnator sum ad

63:2 Wherefore art thou red in thine apparel, and thy gar- saluandum 63:2 quare ergo rubrum est indumentum tuum ments like him that treadeth in the winefat?

63:3 I have trodden the winepress alone; and of the people et uestimenta tua sicut calcantium in torculari 63:3 torcular

there was none with me: for I will tread them in mine anger, and trample them in my fury; and their blood shall be sprinkled upon my garments, and I will stain all my raiment.

calcaui solus et de gentibus non est uir mecum calcaui eos in furore meo et conculcaui eos in ira mea et aspersus est sanguis eorum super uestimenta mea et omnia indumenta

Prophetia Isaiae

mea inquinaui 63:4 dies enim ultionis in corde meo annus 63:4 For the day of vengeance is in mine heart, and the

year of my redeemed is come.

redeeptionis meae uenit 63:5 circumspexi et non erat auxi- 63:5 And I looked, and there was none to help; and I won-

dered that there was none to uphold: therefore mine own arm brought salvation unto me; and my fury, it upheld me.

liator quaesiui et non fuit qui adiuuaret et saluauit mihi brachium meum et indignatio mea ipsa auxiliata est mihi 63:6 63:6 And I will tread down the people in mine anger, and

make them drunk in my fury, and I will bring down their strength to the earth.

et conculcaui populos in furore meo et inebriaui eos in indignatione mea et detraxi in terra uirtutem eorum 63:7 mi- 63:7 I will mention the lovingkindnesses of the LORD,

and the praises of the LORD, according to all that the LORD hath bestowed on us, and the great goodness toward the house of Israel, which he hath bestowed on them according to his mercies, and according to the multitude of his lovingkindnesses.

serationum Domini recordabor laudem Domini super omnibus quae reddidit nobis Dominus et super multitudinem bonorum domui Israhel quae largitus est eis secundum indulgentiam suam et secundum multitudinem misericordiarum suarum 63:8 et dixit uerumtamen populus meus est filii 63:8 For he said, Surely they are my people, children that

will not lie: so he was their Saviour.

non negantes et factus est eis saluator 63:9 in omni tribula- 63:9 In all their affliction he was afflicted, and the angel

of his presence saved them: in his love and in his pity he redeemed them; and he bare them, and carried them all the

tione eorum non est tribulatus et angelus faciei eius saluauit days of old. eos in dilectione sua et in indulgentia sua ipse redeeit eos et portauit eos et leuauit eos cunctis diebus saeculi 63:10 63:10 But they rebelled, and vexed his holy Spirit: there-

fore he was turned to be their enemy, and he fought against them.

ipsi autem ad iracundiam prouocauerunt et adflixerunt spiritum Sancti eius et conuersus est eis in inimicum et ipse debellauit eos 63:11 et recordatus est dierum saeculi Mosi 63:11 Then he remembered the days of old, Moses, and

his people, saying, Where is he that brought them up out of the sea with the shepherd of his flock? where is he that

populi sui ubi est qui eduxit eos de mari cum pastoribus gre- put his holy Spirit within him? gis sui ubi est qui posuit in medio eius spiritum Sancti sui 63:12 qui eduxit ad dexteram Mosen brachio maiestatis suae 63:12 That led them by the right hand of Moses with his

glorious arm, dividing the water before them, to make himself an everlasting name?

qui scidit aquas ante eos ut faceret sibi nomen sempiternum 63:13 qui duxit eos per abyssos quasi equum in deserto non 63:13 That led them through the deep, as an horse in the

wilderness, that they should not stumble?

inpingentem 63:14 quasi animal in campo descendens spi- 63:14 As a beast goeth down into the valley, the Spirit

of the LORD caused him to rest: so didst thou lead thy people, to make thyself a glorious name.

ritus Domini ductor eius fuit sic adduxisti populum tuum ut faceres tibi nomen gloriae 63:15 adtende de caelo et uide 63:15 Look down from heaven, and behold from the habi-

tation of thy holiness and of thy glory: where is thy zeal and thy strength, the sounding of thy bowels and of thy

de habitaculo sancto tuo et gloriae tuae ubi est zelus tuus mercies toward me? are they restrained? et fortitudo tua multitudo uiscerum tuorum et miserationum tuarum super me continuerunt se 63:16 tu enim pater noster 63:16 Doubtless thou art our father, though Abraham be

ignorant of us, and Israel acknowledge us not: thou, O LORD, art our father, our redeemer; thy name is from ev-

et Abraham nesciuit nos et Israhel ignorauit nos tu Domine erlasting. pater noster redeeptor noster a saeculo nomen tuum 63:17 63:17 O LORD, why hast thou made us to err from thy

ways, and hardened our heart from thy fear? Return for thy servants’ sake, the tribes of thine inheritance.

quare errare nos fecisti Domine de uiis tuis indurasti cor nostrum ne timeremus te conuertere propter seruos tuos tribus

Biblia Sacra Vulgata

63:18 The people of thy holiness have possessed it but a hereditatis tuae 63:18 quasi nihilum possederunt populum

little while: our adversaries have trodden down thy sanctuary.

sanctum tuum hostes nostri conculcauerunt sanctificationem

63:19 We are thine: thou never barest rule over them; they tuam 63:19 facti sumus quasi in principio cum non domiwere not called by thy name.

nareris nostri neque inuocaretur nomen tuum super nos

64:1 Oh that thou wouldest rend the heavens, that thou 64:1 utinam disrumperes caelos et descenderes a facie

wouldest come down, that the mountains might flow down at thy presence,

tua montes defluerent 64:2 sicut exustio ignis tabescerent

64:2 As when the melting fire burneth, the fire causeth the waters to boil, to make thy name known to thine adversaries, that the nations may tremble at thy presence!

aquae arderent igni ut notum fieret nomen tuum inimicis

64:3 When thou didst terrible things which we looked not tuis a facie tua gentes turbarentur 64:3 cum feceris mira-

for, thou camest down, the mountains flowed down at thy presence.

bilia non sustinebimus descendisti et a facie tua montes de-

64:4 For since the beginning of the world men have not fluxerunt 64:4 a saeculo non audierunt neque auribus per-

heard, nor perceived by the ear, neither hath the eye seen, O God, beside thee, what he hath prepared for him that

waiteth for him. ceperunt oculus non uidit Deus absque te quae praeparasti 64:5 Thou meetest him that rejoiceth and worketh righ- expectantibus te 64:5 occurristi laetanti et facienti iustitiam

teousness, those that remember thee in thy ways: behold, thou art wroth; for we have sinned: in those is continu-

ance, and we shall be saved. in uiis tuis recordabuntur tui ecce tu iratus es et peccauimus 64:6 But we are all as an unclean thing, and all our righ- in ipsis fuimus semper et saluabimur 64:6 et facti sumus

teousnesses are as filthy rags; and we all do fade as a leaf; and our iniquities, like the wind, have taken us away.

ut inmundus omnes nos quasi pannus menstruatae uniuersae iustitiae nostrae et cecidimus quasi folium uniuersi et iniqui-

64:7 And there is none that calleth upon thy name, that tates nostrae quasi uentus abstulerunt nos 64:7 non est qui

stirreth up himself to take hold of thee: for thou hast hid thy face from us, and hast consumed us, because of our

iniquities. inuocet nomen tuum qui consurgat et teneat te abscondisti

faciem tuam a nobis et adlisisti nos in manu iniquitatis no-

64:8 But now, O LORD, thou art our father; we are the strae 64:8 et nunc Domine pater noster es tu nos uero lutum

clay, and thou our potter; and we all are the work of thy hand.

et fictor noster et opera manuum tuarum omnes nos 64:9

64:9 Be not wroth very sore, O LORD, neither remember iniquity for ever: behold, see, we beseech thee, we are all thy people.

ne irascaris Domine satis et ne ultra memineris iniquitatis

64:10 Thy holy cities are a wilderness, Zion is a wilder- ecce respice populus tuus omnes nos 64:10 ciuitas sancti

ness, Jerusalem a desolation.

tui facta est deserta Sion deserta facta est Hierusalem deso-

64:11 Our holy and our beautiful house, where our fathers lata 64:11 domus sanctificationis nostrae et gloriae nostrae

praised thee, is burned up with fire: and all our pleasant things are laid waste.

ubi laudauerunt te patres nostri facta est in exustionem ig-

64:12 Wilt thou refrain thyself for these things, O LORD? nis et omnia desiderabilia nostra uersa sunt in ruinas 64:12

wilt thou hold thy peace, and afflict us very sore?

numquid super his continebis te Domine tacebis et adfliges nos uehementer

65:1 I am sought of them that asked not for me; I am found 65:1 quaesierunt me qui ante non interrogabant inuenerunt

of them that sought me not: I said, Behold me, behold me, unto a nation that was not called by my name.

qui non quaesierunt me dixi ecce ego ecce ego ad gentem

65:2 I have spread out my hands all the day unto a rebel- quae non uocabat nomen meum 65:2 expandi manus meas lious people, which walketh in a way that was not good, after their own thoughts;

tota die ad populum incredulum qui graditur in uia non bona

Prophetia Isaiae

post cogitationes suas 65:3 populus qui ad iracundiam pro- 65:3 A people that provoketh me to anger continually to

my face; that sacrificeth in gardens, and burneth incense upon altars of brick;

uocat me ante faciem meam semper qui immolant in hortis et sacrificant super lateres 65:4 qui habitant in sepulchris et 65:4 Which remain among the graves, and lodge in the

monuments, which eat swine’s flesh, and broth of abominable things is in their vessels;

in delubris idolorum dormiunt qui comedunt carnem suillam et ius profanum in uasis eorum 65:5 qui dicunt recede a me 65:5 Which say, Stand by thyself, come not near to me;

for I am holier than thou. These are a smoke in my nose, a fire that burneth all the day.

non adpropinques mihi quia inmundus es isti fumus erunt in furore meo ignis ardens tota die 65:6 ecce scriptum est co- 65:6 Behold, it is written before me: I will not keep si-

lence, but will recompense, even recompense into their bosom,

ram me non tacebo sed reddam et retribuam in sinu eorum 65:7 iniquitates uestras et iniquitates patrum uestrorum si- 65:7 Your iniquities, and the iniquities of your fathers to-

gether, saith the LORD, which have burned incense upon the mountains, and blasphemed me upon the hills: there-

mul dicit Dominus qui sacrificauerunt super montes et super fore will I measure their former work into their bosom. colles exprobrauerunt mihi et remetiar opus eorum primum in sinu eorum 65:8 haec dicit Dominus quomodo si inue- 65:8 Thus saith the LORD, As the new wine is found in

the cluster, and one saith, Destroy it not; for a blessing is in it: so will I do for my servants’ sakes, that I may not

niatur granum in botro et dicatur ne dissipes illud quoniam destroy them all. benedictio est sic faciam propter seruos meos ut non dis-

perdam totum 65:9 et educam de Iacob semen et de Iuda 65:9 And I will bring forth a seed out of Jacob, and out of Judah an inheritor of my mountains: and mine elect shall inherit it, and my servants shall dwell there.

possidentem montes meos et hereditabunt eam electi mei et serui mei habitabunt ibi 65:10 et erunt campestria in caulas 65:10 And Sharon shall be a fold of flocks, and the valley

of Achor a place for the herds to lie down in, for my people that have sought me.

gregum et uallis Achor in cubile armentorum populo meo qui requisierunt me 65:11 et uos qui dereliquistis Dominum 65:11 But ye are they that forsake the LORD, that forget

my holy mountain, that prepare a table for that troop, and that furnish the drink offering unto that number.

qui obliti estis montem sanctum meum qui ponitis Fortunae mensam et libatis super eam 65:12 numerabo uos in gladio 65:12 Therefore will I number you to the sword, and ye

shall all bow down to the slaughter: because when I called, ye did not answer; when I spake, ye did not hear; but did evil before mine eyes, and did choose that wherein I delighted not.

et omnes in caede corruetis pro eo quod uocaui et non respondistis locutus sum et non audistis et faciebatis malum in oculis meis et quae nolui elegistis 65:13 propter hoc haec 65:13 Therefore thus saith the Lord GOD, Behold, my ser-

vants shall eat, but ye shall be hungry: behold, my servants shall drink, but ye shall be thirsty: behold, my servants

dicit Dominus Deus ecce serui mei comedent et uos esu- shall rejoice, but ye shall be ashamed: rietis ecce serui mei bibent et uos sitietis 65:14 ecce serui 65:14 Behold, my servants shall sing for joy of heart, but

ye shall cry for sorrow of heart, and shall howl for vexation of spirit.

mei laetabuntur et uos confundemini ecce serui mei laudabunt prae exultatione cordis et uos clamabitis prae dolore cordis et prae contritione spiritus ululabitis 65:15 et dimit- 65:15 And ye shall leave your name for a curse unto my

chosen: for the Lord GOD shall slay thee, and call his servants by another name:

tetis nomen uestrum in iuramentum electis meis et interficiet

te Dominus Deus et seruos suos uocabit nomine alio 65:16 65:16 That he who blesseth himself in the earth shall bless himself in the God of truth; and he that sweareth in the earth shall swear by the God of truth; because the former troubles are forgotten, and because they are hid from mine eyes.

in quo qui benedictus est super terram benedicetur in Deo amen et qui iurat in terra iurabit in Deo amen quia obliuioni

Biblia Sacra Vulgata

traditae sunt angustiae priores et quia absconditae sunt ab

65:17 For, behold, I create new heavens and a new earth: oculis nostris 65:17 ecce enim ego creo caelos nouos et ter-

and the former shall not be remembered, nor come into mind.

ram nouam et non erunt in memoria priora et non ascendent

65:18 But be ye glad and rejoice for ever in that which I super cor 65:18 sed gaudebitis et exultabitis usque in sem-

create: for, behold, I create Jerusalem a rejoicing, and her people a joy.

piternum in his quae ego creo quia ecce ego creo Hierusalem

65:19 And I will rejoice in Jerusalem, and joy in my peo- exultationem et populum eius gaudium 65:19 et exultabo in

ple: and the voice of weeping shall be no more heard in her, nor the voice of crying.

Hierusalem et gaudebo in populo meo et non audietur in eo

65:20 There shall be no more thence an infant of days, nor ultra uox fletus et uox clamoris 65:20 non erit ibi amplius

an old man that hath not filled his days: for the child shall die an hundred years old; but the sinner being an hundred

years old shall be accursed. infans dierum et senex qui non impleat dies suos quoniam

puer centum annorum morietur et peccator centum annorum

65:21 And they shall build houses, and inhabit them; and maledictus erit 65:21 et aedificabunt domos et habitabunt

they shall plant vineyards, and eat the fruit of them. 65:22 They shall not build, and another inhabit; they shall et plantabunt uineas et comedent fructum earum 65:22 non not plant, and another eat: for as the days of a tree are the days of my people, and mine elect shall long enjoy the

work of their hands. aedificabunt et alius habitabit non plantabunt et alius come-

det secundum dies enim ligni erunt dies populi mei et opera

65:23 They shall not labour in vain, nor bring forth for manuum eorum inueterabunt 65:23 electis meis non labora-

trouble; for they are the seed of the blessed of the LORD, and their offspring with them.

bunt frustra neque generabunt in conturbatione quia semen

65:24 And it shall come to pass, that before they call, I benedictorum Domini est et nepotes eorum cum eis 65:24 will answer; and while they are yet speaking, I will hear.

eritque antequam clament ego exaudiam adhuc illis loquen-

65:25 The wolf and the lamb shall feed together, and the tibus ego audiam 65:25 lupus et agnus pascentur simul et

lion shall eat straw like the bullock: and dust shall be the serpent’s meat. They shall not hurt nor destroy in all my

holy mountain, saith the LORD. leo et bos comedent paleas et serpenti puluis panis eius non nocebunt neque occident in omni monte sancto meo dicit Dominus

66:1 Thus saith the LORD, The heaven is my throne, and 66:1 haec dicit Dominus caelum sedis mea et terra sca-

the earth is my footstool: where is the house that ye build unto me? and where is the place of my rest?

billum pedum meorum quae ista domus quam aedificabitis

66:2 For all those things hath mine hand made, and all mihi et quis iste locus quietis meae 66:2 omnia haec manus

those things have been, saith the LORD: but to this man will I look, even to him that is poor and of a contrite spirit,

and trembleth at my word. mea fecit et facta sunt uniuersa ista dicit Dominus ad quem

autem respiciam nisi ad pauperculum et contritum spiritu et

66:3 He that killeth an ox is as if he slew a man; he that trementem sermones meos 66:3 qui immolat bouem quasi

sacrificeth a lamb, as if he cut off a dog’s neck; he that offereth an oblation, as if he offered swine’s blood; he that burneth incense, as if he blessed an idol. Yea, they have chosen their own ways, and their soul delighteth in their abominations.

qui interficiat uirum qui mactat pecus quasi qui excerebret canem qui offert oblationem quasi qui sanguinem suillum offerat qui recordatur turis quasi qui benedicat idolo haec omnia elegerunt in uiis suis et in abominationibus suis an-

66:4 I also will choose their delusions, and will bring their ima eorum delectata est 66:4 unde et ego eligam inlusiones fears upon them; because when I called, none did answer; when I spake, they did not hear: but they did evil before mine eyes, and chose that in which I delighted not.

Prophetia Isaiae

eorum et quae timebant adducam eis quia uocaui et non erat qui responderet locutus sum et non audierunt feceruntque malum in oculis meis et quae nolui elegerunt 66:5 audite 66:5 Hear the word of the LORD, ye that tremble at his

word; Your brethren that hated you, that cast you out for my name’s sake, said, Let the LORD be glorified: but he

uerbum Domini qui tremetis ad uerbum eius dixerunt frat- shall appear to your joy, and they shall be ashamed. res uestri odientes uos et abicientes propter nomen meum glorificetur Dominus et uidebimus in laetitia uestra ipsi autem confundentur 66:6 uox populi de ciuitate uox de tem- 66:6 A voice of noise from the city, a voice from the tem-

ple, a voice of the LORD that rendereth recompence to his enemies.

plo uox Domini reddentis retributionem inimicis suis 66:7

66:7 Before she travailed, she brought forth; before her pain came, she was delivered of a man child.

antequam parturiret peperit antequam ueniret partus eius peperit masculum 66:8 quis audiuit umquam tale et quis ui- 66:8 Who hath heard such a thing? who hath seen such

things? Shall the earth be made to bring forth in one day? or shall a nation be born at once? for as soon as Zion

dit huic simile numquid parturiet terra in die una aut pa- travailed, she brought forth her children. rietur gens simul quia parturiuit et peperit Sion filios suos

66:9 numquid ego qui alios parere facio ipse non pariam di- 66:9 Shall I bring to the birth, and not cause to bring forth? saith the LORD: shall I cause to bring forth, and shut the womb? saith thy God.

cit Dominus si ego qui generationem ceteris tribuo sterilis

ero ait Dominus Deus tuus 66:10 laetamini cum Hierusa- 66:10 Rejoice ye with Jerusalem, and be glad with her, all ye that love her: rejoice for joy with her, all ye that mourn for her:

lem et exultate in ea omnes qui diligitis eam gaudete cum ea gaudio uniuersi qui lugetis super eam 66:11 ut sugatis et 66:11 That ye may suck, and be satisfied with the breasts

of her consolations; that ye may milk out, and be delighted with the abundance of her glory.

repleamini ab ubere consolationis eius ut mulgeatis et deliciis affluatis ab omnimoda gloria eius 66:12 quia haec dicit 66:12 For thus saith the LORD, Behold, I will extend

peace to her like a river, and the glory of the Gentiles like a flowing stream: then shall ye suck, ye shall be borne upon

Dominus ecce ego declinabo super eam quasi fluuium pacis her sides, and be dandled upon her knees. et quasi torrentem inundantem gloriam gentium quam sugetis ad ubera portabimini et super genua blandientur uobis 66:13 quomodo si cui mater blandiatur ita ego consolabor uos 66:13 As one whom his mother comforteth, so will I com-

fort you; and ye shall be comforted in Jerusalem.

et in Hierusalem consolabimini 66:14 uidebitis et gaudebit 66:14 And when ye see this, your heart shall rejoice, and

your bones shall flourish like an herb: and the hand of the LORD shall be known toward his servants, and his

cor uestrum et ossa uestra quasi herba germinabunt et cogno- indignation toward his enemies. scetur manus Domini seruis eius et indignabitur inimicis suis 66:15 quia ecce Dominus in igne ueniet et quasi turbo quadri- 66:15 For, behold, the LORD will come with fire, and with

his chariots like a whirlwind, to render his anger with fury, and his rebuke with flames of fire.

gae eius reddere in indignatione furorem suum et increpationem suam in flamma ignis 66:16 quia in igne Dominus diiu- 66:16 For by fire and by his sword will the LORD plead

with all flesh: and the slain of the LORD shall be many.

dicatur et in gladio suo ad omnem carnem et multiplicabuntur interfecti a Domino 66:17 qui sanctificabantur et mun- 66:17 They that sanctify themselves, and purify them-

selves in the gardens behind one tree in the midst, eating swine’s flesh, and the abomination, and the mouse, shall

dos se putabant in hortis post unam intrinsecus qui comede- be consumed together, saith the LORD. bant carnem suillam et abominationem et murem simul con-

Biblia Sacra Vulgata

66:18 For I know their works and their thoughts: it shall sumentur dicit Dominus 66:18 ego autem opera eorum et come, that I will gather all nations and tongues; and they shall come, and see my glory.

cogitationes eorum uenio ut congregem cum omnibus genti-

66:19 And I will set a sign among them, and I will send bus et linguis et uenient et uidebunt gloriam meam 66:19 et

those that escape of them unto the nations, to Tarshish, Pul, and Lud, that draw the bow, to Tubal, and Javan, to the isles afar off, that have not heard my fame, neither have seen my glory; and they shall declare my glory among the Gentiles.

ponam in eis signum et mittam ex eis qui saluati fuerint ad gentes in mari in Africa in Lydia tenentes sagittam in Italiam et Graeciam ad insulas longe ad eos qui non audierunt de me et non uiderunt gloriam meam et adnuntiabunt gloriam

66:20 And they shall bring all your brethren for an offer- meam gentibus 66:20 et adducent omnes fratres uestros de

ing unto the LORD out of all nations upon horses, and in chariots, and in litters, and upon mules, and upon swift beasts, to my holy mountain Jerusalem, saith the LORD, as the children of Israel bring an offering in a clean vessel into the house of the LORD.

cunctis gentibus donum Domino in equis et in quadrigis et in lecticis et in mulis et in carrucis ad montem sanctum meum Hierusalem dicit Dominus quomodo si inferant filii Israhel

66:21 And I will also take of them for priests and for munus in uase mundo in domum Domini 66:21 et adsumam Levites, saith the LORD.

66:22 For as the new heavens and the new earth, which ex eis in sacerdotes et in Leuitas dicit Dominus 66:22 quia

I will make, shall remain before me, saith the LORD, so shall your seed and your name remain.

sicut caeli noui et terra noua quae ego facio stare coram me dicit Dominus sic stabit semen uestrum et nomen uestrum

66:23 And it shall come to pass, that from one new moon 66:23 et erit mensis ex mense et sabbatum ex sabbato ueniet

to another, and from one sabbath to another, shall all flesh come to worship before me, saith the LORD.

omnis caro ut adoret coram facie mea dicit Dominus 66:24

66:24 And they shall go forth, and look upon the carcases of the men that have transgressed against me: for their worm shall not die, neither shall their fire be quenched; and they shall be an abhorring unto all flesh.

et egredientur et uidebunt cadauera uirorum qui praeuaricati sunt in me uermis eorum non morietur et ignis eorum non extinguetur et erunt usque ad satietatem uisionis omni carni

Prophetia Ieremiae

1:1 uerba Hieremiae filii Helciae de sacerdotibus qui fue- 1:1 The words of Jeremiah the son of Hilkiah, of the

priests that were in Anathoth in the land of Benjamin:

runt in Anathoth in terra Beniamin 1:2 quod factum est uer- 1:2 To whom the word of the LORD came in the days of

Josiah the son of Amon king of Judah, in the thirteenth year of his reign.

bum Domini ad eum in diebus Iosiae filii Amon regis Iuda in tertiodecimo anno regni eius 1:3 et factum est in diebus 1:3 It came also in the days of Jehoiakim the son of

Josiah king of Judah, unto the end of the eleventh year of Zedekiah the son of Josiah king of Judah, unto the car-

Ioachim filii Iosiae regis Iuda usque ad consummationem rying away of Jerusalem captive in the fifth month. undecimi anni Sedeciae filii Iosiae regis Iuda usque ad transmigrationem Hierusalem in mense quinto 1:4 et factum est 1:4 Then the word of the LORD came unto me, saying, uerbum Domini ad me dicens 1:5 priusquam te formarem 1:5 Before I formed thee in the belly I knew thee; and

before thou camest forth out of the womb I sanctified thee, and I ordained thee a prophet unto the nations.

in utero noui te et antequam exires de uulua sanctificaui te prophetam gentibus dedi te 1:6 et dixi a a a Domine Deus 1:6 Then said I, Ah, Lord GOD! behold, I cannot speak:

for I am a child.

ecce nescio loqui quia puer ego sum 1:7 et dixit Dominus 1:7 But the LORD said unto me, Say not, I am a child: for

thou shalt go to all that I shall send thee, and whatsoever I command thee thou shalt speak.

ad me noli dicere puer sum quoniam ad omnia quae mittam te ibis et uniuersa quaecumque mandauero tibi loqueris 1:8 1:8 Be not afraid of their faces: for I am with thee to de-

liver thee, saith the LORD.

ne timeas a facie eorum quia tecum ego sum ut eruam te dicit Dominus 1:9 et misit Dominus manum suam et tetigit 1:9 Then the LORD put forth his hand, and touched my

mouth. And the LORD said unto me, Behold, I have put my words in thy mouth.

os meum et dixit Dominus ad me ecce dedi uerba mea in ore tuo 1:10 ecce constitui te hodie super gentes et super 1:10 See, I have this day set thee over the nations and over

the kingdoms, to root out, and to pull down, and to destroy, and to throw down, to build, and to plant.

regna ut euellas et destruas et disperdas et dissipes et aedifices et plantes 1:11 et factum est uerbum Domini ad me 1:11 Moreover the word of the LORD came unto me, say-

ing, Jeremiah, what seest thou? And I said, I see a rod of an almond tree.

dicens quid tu uides Hieremia et dixi uirgam uigilantem ego uideo 1:12 et dixit Dominus ad me bene uidisti quia uigi- 1:12 Then said the LORD unto me, Thou hast well seen:

for I will hasten my word to perform it.

labo ego super uerbo meo ut faciam illud 1:13 et factum est 1:13 And the word of the LORD came unto me the second

time, saying, What seest thou? And I said, I see a seething pot; and the face thereof is toward the north.

uerbum Domini secundo ad me dicens quid tu uides et dixi ollam succensam ego uideo et faciem eius a facie aquilonis 1:14 et dixit Dominus ad me ab aquilone pandetur malum 1:14 Then the LORD said unto me, Out of the north an

evil shall break forth upon all the inhabitants of the land.

super omnes habitatores terrae 1:15 quia ecce ego conuo- 1:15 For, lo, I will call all the families of the kingdoms of the north, saith the LORD; and they shall come, and they shall set every one his throne at the entering of the gates of Jerusalem, and against all the walls thereof round about, and against all the cities of Judah.

cabo omnes cognationes regnorum aquilonis ait Dominus et uenient et ponent unusquisque solium suum in introitu portarum Hierusalem et super omnes muros eius in circuitu et super uniuersas urbes Iuda 1:16 et loquar iudicia mea cum 1:16 And I will utter my judgments against them touch-

ing all their wickedness, who have forsaken me, and have burned incense unto other gods, and worshipped the works

eis super omni malitia eorum qui dereliquerunt me et libaue- of their own hands.

runt diis alienis et adorauerunt opus manuum suarum 1:17 1:17 Thou therefore gird up thy loins, and arise, and speak unto them all that I command thee: be not dismayed at their faces, lest I confound thee before them.

tu ergo accinge lumbos tuos et surge et loquere ad eos om-

nia quae ego praecipio tibi ne formides a facie eorum nec

1:18 For, behold, I have made thee this day a defenced city, enim timere te faciam uultum eorum 1:18 ego quippe deand an iron pillar, and brasen walls against the whole land, against the kings of Judah, against the princes thereof, against the priests thereof, and against the people of the land.

di te hodie in ciuitatem munitam et in columnam ferream et in murum aereum super omnem terram regibus Iuda princi-

1:19 And they shall fight against thee; but they shall not pibus eius et sacerdotibus et populo terrae 1:19 et bellabunt

prevail against thee; for I am with thee, saith the LORD, to deliver thee.

aduersum te et non praeualebunt quia tecum ego sum ait Dominus ut liberem te

2:1 Moreover the word of the LORD came to me, saying, 2:1 et factum est uerbum Domini ad me dicens 2:2 uade

2:2 Go and cry in the ears of Jerusalem, saying, Thus saith the LORD; I remember thee, the kindness of thy youth, the love of thine espousals, when thou wentest after me in the wilderness, in a land that was not sown.

et clama in auribus Hierusalem dicens haec dicit Dominus recordatus sum tui miserans adulescentiam tuam et caritatem disponsationis tuae quando secuta me es in deserto in terra

2:3 Israel was holiness unto the LORD, and the firstfruits quae non seminatur 2:3 sanctus Israhel Domino primitiae of his increase: all that devour him shall offend; evil shall come upon them, saith the LORD.

frugum eius omnes qui deuorant eum delinquunt mala ue-

2:4 Hear ye the word of the LORD, O house of Jacob, and nient super eos dicit Dominus 2:4 audite uerbum Domini

all the families of the house of Israel:

2:5 Thus saith the LORD, What iniquity have your fathers domus Iacob et omnes cognationes domus Israhel 2:5 haec found in me, that they are gone far from me, and have walked after vanity, and are become vain?

dicit Dominus quid inuenerunt patres uestri in me iniquitatis quia elongauerunt a me et ambulauerunt post uanitatem

2:6 Neither said they, Where is the LORD that brought et uani facti sunt 2:6 et non dixerunt ubi est Dominus qui

us up out of the land of Egypt, that led us through the wilderness, through a land of deserts and of pits, through a land of drought, and of the shadow of death, through a land that no man passed through, and where no man dwelt?

ascendere nos fecit de terra Aegypti qui transduxit nos per desertum per terram inhabitabilem et inuiam per terram sitis et imaginem mortis per terram in qua non ambulauit uir ne-

2:7 And I brought you into a plentiful country, to eat the que habitauit homo 2:7 et induxi uos in terram Carmeli ut fruit thereof and the goodness thereof; but when ye entered, ye defiled my land, and made mine heritage an

abomination. comederetis fructum eius et optima illius et ingressi contaminastis terram meam et hereditatem meam posuistis in abo-

2:8 The priests said not, Where is the LORD? and they minationem 2:8 sacerdotes non dixerunt ubi est Dominus et

that handle the law knew me not: the pastors also transgressed against me, and the prophets prophesied by Baal,

and walked after things that do not profit. tenentes legem nescierunt me et pastores praeuaricati sunt in me et prophetae prophetauerunt in Baal et idola secuti sunt

2:9 Wherefore I will yet plead with you, saith the LORD, 2:9 propterea adhuc iudicio contendam uobiscum ait Domi-

and with your children’s children will I plead.

2:10 For pass over the isles of Chittim, and see; and send nus et cum filiis uestris disceptabo 2:10 transite ad insulas

unto Kedar, and consider diligently, and see if there be such a thing.

Cetthim et uidete et in Cedar mittite et considerate uehemen-

2:11 Hath a nation changed their gods, which are yet no ter et uidete si factum est huiuscemodi 2:11 si mutauit gens

gods? but my people have changed their glory for that which doth not profit.

deos et certe ipsi non sunt dii populus uero meus mutauit

2:12 Be astonished, O ye heavens, at this, and be horribly Gloriam suam in idolum 2:12 obstupescite caeli super hoc afraid, be ye very desolate, saith the LORD.

Prophetia Ieremiae

et portae eius desolamini uehementer dicit Dominus 2:13 2:13 For my people have committed two evils; they have

forsaken me the fountain of living waters, and hewed them out cisterns, broken cisterns, that can hold no water.

duo enim mala fecit populus meus me dereliquerunt fontem aquae uiuae ut foderent sibi cisternas cisternas dissipatas quae continere non ualent aquas 2:14 numquid seruus 2:14 Is Israel a servant? is he a homeborn slave? why is

he spoiled?

est Israhel aut uernaculus quare ergo est factus in praedam 2:15 super eum rugierunt leones et dederunt uocem suam po- 2:15 The young lions roared upon him, and yelled, and

they made his land waste: his cities are burned without inhabitant.

suerunt terram eius in solitudinem ciuitates eius exustae sunt et non est qui habitet in eis 2:16 filii quoque Memfeos et 2:16 Also the children of Noph and Tahapanes have bro-

ken the crown of thy head.

Tafnes constuprauerunt te usque ad uerticem 2:17 numquid 2:17 Hast thou not procured this unto thyself, in that thou

hast forsaken the LORD thy God, when he led thee by the way?

non istud factum est tibi quia dereliquisti Dominum Deum tuum eo tempore quo ducebat te per uiam 2:18 et nunc quid 2:18 And now what hast thou to do in the way of Egypt,

to drink the waters of Sihor? or what hast thou to do in the way of Assyria, to drink the waters of the river?

tibi uis in uia Aegypti ut bibas aquam turbidam et quid tibi cum uia Assyriorum ut bibas aquam Fluminis 2:19 arguet 2:19 Thine own wickedness shall correct thee, and thy

backslidings shall reprove thee: know therefore and see that it is an evil thing and bitter, that thou hast forsaken the LORD thy God, and that my fear is not in thee, saith the Lord GOD of hosts.

te malitia tua et auersio tua increpabit te scito et uide quia malum et amarum est reliquisse te Dominum Deum tuum et non esse timorem mei apud te dicit Dominus Deus exercituum 2:20 a saeculo confregisti iugum meum rupisti uin- 2:20 For of old time I have broken thy yoke, and burst thy

bands; and thou saidst, I will not transgress; when upon every high hill and under every green tree thou wanderest,

cula mea et dixisti non seruiam in omni enim colle sublimi playing the harlot. et sub omni ligno frondoso tu prosternebaris meretrix 2:21 2:21 Yet I had planted thee a noble vine, wholly a right

seed: how then art thou turned into the degenerate plant of a strange vine unto me?

ego autem plantaui te uineam electam omne semen uerum quomodo ergo conuersa es in prauum uinea aliena 2:22 si 2:22 For though thou wash thee with nitre, and take thee

much soap, yet thine iniquity is marked before me, saith the Lord GOD.

laueris te nitro et multiplicaueris tibi herbam borith maculata es in iniquitate tua coram me dicit Dominus Deus 2:23 2:23 How canst thou say, I am not polluted, I have not

gone after Baalim? see thy way in the valley, know what thou hast done: thou art a swift dromedary traversing her

quomodo dicis non sum polluta post Baalim non ambulaui ways; uide uias tuas in conualle scito quid feceris cursor leuis explicans uias tuas 2:24 onager adsuetus in solitudine in desi- 2:24 A wild ass used to the wilderness, that snuffeth up

the wind at her pleasure; in her occasion who can turn her away? all they that seek her will not weary themselves; in

derio animae suae adtraxit uentum amoris sui nullus auertet her month they shall find her. eam omnes qui quaerunt eam non deficient in menstruis eius inuenient eam 2:25 prohibe pedem tuum a nuditate et guttur 2:25 Withhold thy foot from being unshod, and thy throat

from thirst: but thou saidst, There is no hope: no; for I have loved strangers, and after them will I go.

tuum a siti et dixisti desperaui nequaquam faciam adamaui quippe alienos et post eos ambulabo 2:26 quomodo confun- 2:26 As the thief is ashamed when he is found, so is the

house of Israel ashamed; they, their kings, their princes, and their priests, and their prophets.

ditur fur quando deprehenditur sic confusi sunt domus Israhel ipsi et reges eorum principes et sacerdotes et prophetae

Biblia Sacra Vulgata

2:27 Saying to a stock, Thou art my father; and to a stone, eorum 2:27 dicentes ligno pater meus es tu et lapidi tu me Thou hast brought me forth: for they have turned their back unto me, and not their face: but in the time of their

trouble they will say, Arise, and save us. genuisti uerterunt ad me tergum et non faciem et in tempore 2:28 But where are thy gods that thou hast made thee? let adflictionis suae dicent surge et libera nos 2:28 ubi sunt dii

them arise, if they can save thee in the time of thy trouble: for according to the number of thy cities are thy gods, O

Judah. tui quos fecisti tibi surgant et liberent te in tempore adflic-

tionis tuae secundum numerum quippe ciuitatum tuarum er-

2:29 Wherefore will ye plead with me? ye all have trans- ant dii tui Iuda 2:29 quid uultis mecum iudicio contendere

gressed against me, saith the LORD.

2:30 In vain have I smitten your children; they received no omnes dereliquistis me dicit Dominus 2:30 frustra percussi

correction: your own sword hath devoured your prophets, like a destroying lion.

filios uestros disciplinam non receperunt deuorauit gladius

2:31 O generation, see ye the word of the LORD. Have I uester prophetas uestros quasi leo uastator 2:31 generatio

been a wilderness unto Israel? a land of darkness? wherefore say my people, We are lords; we will come no more

unto thee? uestra uidete uerbum Domini numquid solitudo factus sum

Israheli aut terra serotina quare ergo dixit populus meus re-

2:32 Can a maid forget her ornaments, or a bride her attire? cessimus non ueniemus ultra ad te 2:32 numquid obliuisciyet my people have forgotten me days without number.

tur uirgo ornamenti sui sponsa fasciae pectoralis suae popu-

2:33 Why trimmest thou thy way to seek love? therefore lus uero meus oblitus est mei diebus innumeris 2:33 quid

hast thou also taught the wicked ones thy ways.

niteris bonam ostendere uiam tuam ad quaerendam dilectio-

2:34 Also in thy skirts is found the blood of the souls of nem quae insuper et malitias tuas docuisti uias tuas 2:34 et

the poor innocents: I have not found it by secret search, but upon all these.

in alis tuis inuentus est sanguis animarum pauperum et innocentium non in fossis inueni eos sed in omnibus quae supra

2:35 Yet thou sayest, Because I am innocent, surely his memoraui 2:35 et dixisti absque peccato et innocens ego

anger shall turn from me. Behold, I will plead with thee, because thou sayest, I have not sinned.

sum et propterea auertatur furor tuus a me ecce ego iudicio

2:36 Why gaddest thou about so much to change thy way? contendam tecum eo quod dixeris non peccaui 2:36 quam

thou also shalt be ashamed of Egypt, as thou wast ashamed of Assyria.

uilis es facta nimis iterans uias tuas et ab Aegypto confunde-

2:37 Yea, thou shalt go forth from him, and thine hands ris sicut confusa es ab Assur 2:37 nam et ab ista egredieris

upon thine head: for the LORD hath rejected thy confidences, and thou shalt not prosper in them.

et manus tuae erunt super caput tuum quoniam obtriuit Dominus confidentiam tuam et nihil habebis prosperum

3:1 They say, If a man put away his wife, and she go from 3:1 uulgo dicitur si dimiserit uir uxorem suam et recedens

him, and become another man’s, shall he return unto her again? shall not that land be greatly polluted? but thou hast played the harlot with many lovers; yet return again to me, saith the LORD.

ab eo duxerit uirum alterum numquid reuertetur ad eam ultra numquid non polluta et contaminata erit mulier illa tu autem fornicata es cum amatoribus multis tamen reuertere ad me

3:2 Lift up thine eyes unto the high places, and see where dicit Dominus 3:2 leua oculos tuos in directum et uide ubi

thou hast not been lien with. In the ways hast thou sat for them, as the Arabian in the wilderness; and thou hast polluted the land with thy whoredoms and with thy wickedness.

non prostrata sis in uiis sedebas expectans eos quasi latro in solitudine et polluisti terram in fornicationibus tuis et in ma-

3:3 Therefore the showers have been withholden, and litiis tuis 3:3 quam ob rem prohibitae sunt stillae pluuiarum

there hath been no latter rain; and thou hadst a whore’s forehead, thou refusedst to be ashamed.

Prophetia Ieremiae

et serotinus imber non fuit frons mulieris meretricis facta est tibi noluisti erubescere 3:4 ergo saltim amodo uoca me pa- 3:4 Wilt thou not from this time cry unto me, My father,

thou art the guide of my youth?

ter meus dux uirginitatis meae tu es 3:5 numquid irasceris 3:5 Will he reserve his anger for ever? will he keep it to

the end? Behold, thou hast spoken and done evil things as thou couldest.

in perpetuum aut perseuerabis in finem ecce locuta es et fecisti mala et potuisti 3:6 et dixit Dominus ad me in diebus 3:6 The LORD said also unto me in the days of Josiah

the king, Hast thou seen that which backsliding Israel hath done? she is gone up upon every high mountain and under

Iosiae regis numquid uidisti quae fecerit auersatrix Israhel every green tree, and there hath played the harlot. abiit sibimet super omnem montem excelsum et sub omne lignum frondosum et fornicata est ibi 3:7 et dixi cum fecis- 3:7 And I said after she had done all these things, Turn

thou unto me. But she returned not. And her treacherous sister Judah saw it.

set haec omnia ad me conuertere et non est reuersa et uidit praeuaricatrix soror eius Iuda 3:8 quia pro eo quod moe- 3:8 And I saw, when for all the causes whereby back-

sliding Israel committed adultery I had put her away, and given her a bill of divorce; yet her treacherous sister Judah

chata esset auersatrix Israhel dimisissem eam et dedissem ei feared not, but went and played the harlot also. libellum repudii et non timuit praeuaricatrix Iuda soror eius sed abiit et fornicata est etiam ipsa 3:9 et facilitate fornica- 3:9 And it came to pass through the lightness of her whore-

dom, that she defiled the land, and committed adultery with stones and with stocks.

tionis suae contaminauit terram et moechata est cum lapide

et cum ligno 3:10 et in omnibus his non est reuersa ad me 3:10 And yet for all this her treacherous sister Judah hath not turned unto me with her whole heart, but feignedly, saith the LORD.

praeuaricatrix soror eius Iuda in toto corde suo sed in mendacio ait Dominus 3:11 et dixit Dominus ad me iustificauit 3:11 And the LORD said unto me, The backsliding Israel

hath justified herself more than treacherous Judah.

animam suam auersatrix Israhel conparatione praeuaricatricis Iuda 3:12 uade et clama sermones istos contra aquilo- 3:12 Go and proclaim these words toward the north, and

say, Return, thou backsliding Israel, saith the LORD; and I will not cause mine anger to fall upon you: for I am merciful, saith the LORD, and I will not keep anger for ever.

nem et dices reuertere auersatrix Israhel ait Dominus et non auertam faciem meam a uobis quia sanctus ego sum dicit

Dominus et non irascar in perpetuum 3:13 tamen scito in- 3:13 Only acknowledge thine iniquity, that thou hast transgressed against the LORD thy God, and hast scattered thy ways to the strangers under every green tree, and ye have

iquitatem tuam quia in Dominum Deum tuum praeuaricata not obeyed my voice, saith the LORD. es et dispersisti uias tuas alienis sub omni ligno frondoso et uocem meam non audisti ait Dominus 3:14 conuertimini fi- 3:14 Turn, O backsliding children, saith the LORD; for I

am married unto you: and I will take you one of a city, and two of a family, and I will bring you to Zion:

lii reuertentes dicit Dominus quia ego uir uester et adsumam uos unum de ciuitate et duos de cognatione et introducam

uos in Sion 3:15 et dabo uobis pastores iuxta cor meum et 3:15 And I will give you pastors according to mine heart,

which shall feed you with knowledge and understanding.

pascent uos scientia et doctrina 3:16 cumque multiplicati 3:16 And it shall come to pass, when ye be multiplied and

increased in the land, in those days, saith the LORD, they shall say no more, The ark of the covenant of the LORD: neither shall it come to mind: neither shall they remember it; neither shall they visit it; neither shall that be done any more.

fueritis et creueritis in terra in diebus illis ait Dominus non dicent ultra arca testamenti Domini neque ascendet super cor neque recordabuntur illius nec uisitabitur nec fiet ultra 3:17 3:17 At that time they shall call Jerusalem the throne of

the LORD; and all the nations shall be gathered unto it, to the name of the LORD, to Jerusalem: neither shall they

in tempore illo uocabunt Hierusalem solium Domini et con- walk any more after the imagination of their evil heart.

Biblia Sacra Vulgata

gregabuntur ad eam omnes gentes in nomine Domini in Hierusalem et non ambulabunt post prauitatem cordis sui pes-

3:18 In those days the house of Judah shall walk with the simi 3:18 in diebus illis ibit domus Iuda ad domum Israhel

house of Israel, and they shall come together out of the land of the north to the land that I have given for an inher-

itance unto your fathers. et uenient simul de terra aquilonis ad terram quam dedi patri-3:19 But I said, How shall I put thee among the children, bus uestris 3:19 ego autem dixi quomodo ponam te in filiis

and give thee a pleasant land, a goodly heritage of the hosts of nations? and I said, Thou shalt call me, My fa-

ther; and shalt not turn away from me. et tribuam tibi terram desiderabilem hereditatem praeclaram

exercituum gentium et dixi patrem uocabis me et post me

3:20 Surely as a wife treacherously departeth from her ingredi non cessabis 3:20 sed quomodo si contemnat muhusband, so have ye dealt treacherously with me, O house of Israel, saith the LORD.

lier amatorem suum sic contempsit me domus Israhel dicit

3:21 A voice was heard upon the high places, weeping Dominus 3:21 uox in uiis audita est ploratus et ululatus fi-

and supplications of the children of Israel: for they have perverted their way, and they have forgotten the LORD

their God. liorum Israhel quoniam iniquam fecerunt uiam suam obliti

3:22 Return, ye backsliding children, and I will heal your sunt Domini Dei sui 3:22 conuertimini filii reuertentes et backslidings. Behold, we come unto thee; for thou art the LORD our God.

sanabo auersiones uestras ecce nos uenimus ad te tu enim

3:23 Truly in vain is salvation hoped for from the hills, and es Dominus Deus noster 3:23 uere mendaces erant colles from the multitude of mountains: truly in the LORD our God is the salvation of Israel.

multitudo montium uere in Domino Deo nostro salus Israhel

3:24 For shame hath devoured the labour of our fathers 3:24 confusio comedit laborem patrum nostrorum ab adules-

from our youth; their flocks and their herds, their sons and their daughters.

centia nostra greges eorum et armenta eorum filios eorum et

3:25 We lie down in our shame, and our confusion cov- filias eorum 3:25 dormiemus in confusione nostra et operiet

ereth us: for we have sinned against the LORD our God, we and our fathers, from our youth even unto this day, and

have not obeyed the voice of the LORD our God. nos ignominia nostra quoniam Domino Deo nostro peccauimus nos et patres nostri ab adulescentia nostra usque ad hanc diem et non audiuimus uocem Domini Dei nostri

4:1 If thou wilt return, O Israel, saith the LORD, return 4:1 si conuerteris Israhel ait Dominus ad me conuertere unto me: and if thou wilt put away thine abominations out of my sight, then shalt thou not remove.

si abstuleris offendicula tua a facie mea non commoueberis

4:2 And thou shalt swear, The LORD liveth, in truth, in 4:2 et iurabis uiuit Dominus in ueritate et in iudicio et in

judgment, and in righteousness; and the nations shall bless themselves in him, and in him shall they glory.

iustitia et benedicent eum gentes ipsumque laudabunt 4:3

4:3 For thus saith the LORD to the men of Judah and Jerusalem, Break up your fallow ground, and sow not among thorns.

haec enim dicit Dominus uiro Iuda et Hierusalem nouate uo-

4:4 Circumcise yourselves to the LORD, and take away bis nouale et nolite serere super spinas 4:4 circumcidimini

the foreskins of your heart, ye men of Judah and inhabitants of Jerusalem: lest my fury come forth like fire, and burn that none can quench it, because of the evil of your doings.

Domino et auferte praeputia cordium uestrorum uir Iuda et habitatores Hierusalem ne forte egrediatur ut ignis indignatio mea et succendatur et non sit qui extinguat propter ma-

4:5 Declare ye in Judah, and publish in Jerusalem; and say, litiam cogitationum uestrarum 4:5 adnuntiate in Iuda et in

Blow ye the trumpet in the land: cry, gather together, and say, Assemble yourselves, and let us go into the defenced

cities. Hierusalem auditum facite loquimini et canite tuba in terra clamate fortiter dicite congregamini et ingrediamur ciuita-

Prophetia Ieremiae

tes munitas 4:6 leuate signum in Sion confortamini nolite 4:6 Set up the standard toward Zion: retire, stay not: for I

will bring evil from the north, and a great destruction.

stare quia malum ego adduco ab aquilone et contritionem magnam 4:7 ascendit leo de cubili suo et praedo gentium se 4:7 The lion is come up from his thicket, and the destroyer

of the Gentiles is on his way; he is gone forth from his place to make thy land desolate; and thy cities shall be

leuauit egressus est de loco suo ut ponat terram tuam in deso- laid waste, without an inhabitant. lationem ciuitates tuae uastabuntur remanentes absque habitatore 4:8 super hoc accingite uos ciliciis plangite et ululate 4:8 For this gird you with sackcloth, lament and howl: for

the fierce anger of the LORD is not turned back from us.

quia non est auersa ira furoris Domini a nobis 4:9 et erit in 4:9 And it shall come to pass at that day, saith the LORD,

that the heart of the king shall perish, and the heart of the princes; and the priests shall be astonished, and the

die illa dicit Dominus peribit cor regis et cor principum et prophets shall wonder. obstupescent sacerdotes et prophetae consternabuntur 4:10 4:10 Then said I, Ah, Lord GOD! surely thou hast greatly

deceived this people and Jerusalem, saying, Ye shall have peace; whereas the sword reacheth unto the soul.

et dixi heu heu heu Domine Deus ergone decepisti populum istum et Hierusalem dicens pax erit uobis et ecce peruenit gladius usque ad animam 4:11 in tempore illo dicetur po- 4:11 At that time shall it be said to this people and to

Jerusalem, A dry wind of the high places in the wilderness toward the daughter of my people, not to fan, nor to

pulo huic et Hierusalem uentus urens in uiis quae sunt in cleanse, deserto uiae filiae populi mei non ad uentilandum et ad purgandum 4:12 spiritus plenus ex his ueniet mihi et nunc ego 4:12 Even a full wind from those places shall come unto

me: now also will I give sentence against them.

sed loquar iudicia mea cum eis 4:13 ecce quasi nubes as- 4:13 Behold, he shall come up as clouds, and his chariots shall be as a whirlwind: his horses are swifter than eagles. Woe unto us! for we are spoiled.

cendet et quasi tempestas currus eius uelociores aquilis equi

illius uae nobis quoniam uastati sumus 4:14 laua a mali- 4:14 O Jerusalem, wash thine heart from wickedness, that thou mayest be saved. How long shall thy vain thoughts lodge within thee?

tia cor tuum Hierusalem ut salua fias usquequo morabuntur in te cogitationes noxiae 4:15 uox enim adnuntiantis a 4:15 For a voice declareth from Dan, and publisheth af-

fliction from mount Ephraim.

Dan et notum facientis idolum de monte Ephraim 4:16 con- 4:16 Make ye mention to the nations; behold, publish

against Jerusalem, that watchers come from a far country, and give out their voice against the cities of Judah.

citate gentes ecce auditum est in Hierusalem custodes uenire de terra longinqua et dare super ciuitates Iuda uocem suam 4:17 quasi custodes agrorum facti sunt super eam in 4:17 As keepers of a field, are they against her round

about; because she hath been rebellious against me, saith the LORD.

gyro quia me ad iracundiam prouocauit ait Dominus 4:18

4:18 Thy way and thy doings have procured these things unto thee; this is thy wickedness, because it is bitter, because it reacheth unto thine heart.

uiae tuae et cogitationes tuae fecerunt haec tibi ista malitia tua quia amara quia tetigit cor tuum 4:19 uentrem meum 4:19 My bowels, my bowels! I am pained at my very heart;

my heart maketh a noise in me; I cannot hold my peace, because thou hast heard, O my soul, the sound of the trum-

uentrem meum doleo sensus cordis mei turbati sunt in me pet, the alarm of war. non tacebo quoniam uocem bucinae audiuit anima mea clamorem proelii 4:20 contritio super contritionem uocata est 4:20 Destruction upon destruction is cried; for the whole

land is spoiled: suddenly are my tents spoiled, and my curtains in a moment.

et uastata est omnis terra repente uastata sunt tabernacula

mea subito pelles meae 4:21 usquequo uidebo fugientem 4:21 How long shall I see the standard, and hear the sound

of the trumpet?

audiam uocem bucinae 4:22 quia stultus populus meus me 4:22 For my people is foolish, they have not known me;

they are sottish children, and they have none understanding: they are wise to do evil, but to do good they have no knowledge.

Biblia Sacra Vulgata

non cognouit filii insipientes sunt et uecordes sapientes sunt

4:23 I beheld the earth, and, lo, it was without form, and ut faciant mala bene autem facere nescierunt 4:23 aspexi

void; and the heavens, and they had no light.

terram et ecce uacua erat et nihili et caelos et non erat lux

4:24 I beheld the mountains, and, lo, they trembled, and in eis 4:24 uidi montes et ecce mouebantur et omnes colall the hills moved lightly. 4:25 I beheld, and, lo, there was no man, and all the birds les conturbati sunt 4:25 intuitus sum et non erat homo et of the heavens were fled. 4:26 I beheld, and, lo, the fruitful place was a wilderness, omne uolatile caeli recessit 4:26 aspexi et ecce Carmelus and all the cities thereof were broken down at the presence of the LORD, and by his fierce anger.

desertus et omnes urbes eius destructae sunt a facie Domini

4:27 For thus hath the LORD said, The whole land shall et a facie irae furoris eius 4:27 haec enim dicit Dominus

be desolate; yet will I not make a full end.

deserta erit omnis terra sed tamen consummationem non fa-

4:28 For this shall the earth mourn, and the heavens above ciam 4:28 lugebit terra et maerebunt caeli desuper eo quod

be black; because I have spoken it, I have purposed it, and will not repent, neither will I turn back from it.

locutus sum cogitaui et non paenituit me nec auersus sum

4:29 The whole city shall flee for the noise of the horse- ab eo 4:29 a uoce equitis et mittentis sagittam fugit omnis

men and bowmen; they shall go into thickets, and climb up upon the rocks: every city shall be forsaken, and not a

man dwell therein. ciuitas ingressi sunt ardua et ascenderunt rupes uniuersae ur-4:30 And when thou art spoiled, what wilt thou do? bes derelictae sunt et non habitat in eis homo 4:30 tu autem

Though thou clothest thyself with crimson, though thou deckest thee with ornaments of gold, though thou rentest thy face with painting, in vain shalt thou make thyself fair; thy lovers will despise thee, they will seek thy life.

uastata quid facies cum uestieris te coccino cum ornata fueris monili aureo et pinxeris stibio oculos tuos frustra conponeris contempserunt te amatores tui animam tuam quaerent

4:31 For I have heard a voice as of a woman in travail, and 4:31 uocem enim quasi parturientis audiui angustias ut puer-

the anguish as of her that bringeth forth her first child, the voice of the daughter of Zion, that bewaileth herself, that spreadeth her hands, saying, Woe is me now! for my soul is wearied because of murderers.

perae uox filiae Sion intermorientis expandentisque manus suas uae mihi quia defecit anima mea propter interfectos

5:1 Run ye to and fro through the streets of Jerusalem, and 5:1 circuite uias Hierusalem et aspicite et considerate et

see now, and know, and seek in the broad places thereof, if ye can find a man, if there be any that executeth judgment,

that seeketh the truth; and I will pardon it. quaerite in plateis eius an inueniatis uirum facientem iudi-5:2 And though they say, The LORD liveth; surely they cium et quaerentem fidem et propitius ero eius 5:2 quod si

swear falsely.

5:3 O LORD, are not thine eyes upon the truth? thou hast etiam uiuit Dominus dixerint et hoc falso iurabunt 5:3 Do-

stricken them, but they have not grieved; thou hast consumed them, but they have refused to receive correction: they have made their faces harder than a rock; they have refused to return.

mine oculi tui respiciunt fidem percussisti eos et non doluerunt adtriuisti eos et rennuerunt accipere disciplinam indu-

5:4 Therefore I said, Surely these are poor; they are fool- rauerunt facies suas super petram noluerunt reuerti 5:4 ego

ish: for they know not the way of the LORD, nor the judgment of their God.

autem dixi forsitan pauperes sunt et stulti ignorantes uiam

5:5 I will get me unto the great men, and will speak unto Domini iudicium Dei sui 5:5 ibo igitur ad optimates et lo-

them; for they have known the way of the LORD, and the judgment of their God: but these have altogether broken

the yoke, and burst the bonds. quar eis ipsi enim cognouerunt uiam Domini iudicium Dei

sui et ecce magis hii simul confregerunt iugum ruperunt uin-

5:6 Wherefore a lion out of the forest shall slay them, and cula 5:6 idcirco percussit eos leo de silua lupus ad uespe-

a wolf of the evenings shall spoil them, a leopard shall watch over their cities: every one that goeth out thence shall be torn in pieces: because their transgressions are many, and their backslidings are increased.

ram uastauit eos pardus uigilans super ciuitates eorum om-

Prophetia Ieremiae

nis qui egressus fuerit ex eis capietur quia multiplicatae sunt praeuaricationes eorum confortatae sunt auersiones eorum 5:7 super quo propitius tibi esse potero filii tui dereliquerunt 5:7 How shall I pardon thee for this? thy children have

forsaken me, and sworn by them that are no gods: when I had fed them to the full, they then committed adultery, and

me et iurant in his qui non sunt dii saturaui eos et moechati assembled themselves by troops in the harlots’ houses. sunt et in domo meretricis luxuriabantur 5:8 equi amatores 5:8 They were as fed horses in the morning: every one

neighed after his neighbour’s wife.

et admissarii facti sunt unusquisque ad uxorem proximi sui hinniebat 5:9 numquid super his non uisitabo dicit Domi- 5:9 Shall I not visit for these things? saith the LORD: and

shall not my soul be avenged on such a nation as this?

nus et in gente tali non ulciscetur anima mea 5:10 ascen- 5:10 Go ye up upon her walls, and destroy; but make not

a full end: take away her battlements; for they are not the LORD’s.

dite muros eius et dissipate consummationem autem nolite facere auferte propagines eius quia non sunt Domini 5:11 5:11 For the house of Israel and the house of Judah have

dealt very treacherously against me, saith the LORD.

praeuaricatione enim praeuaricata est in me domus Israhel et domus Iuda ait Dominus 5:12 negauerunt Dominum et 5:12 They have belied the LORD, and said, It is not he;

neither shall evil come upon us; neither shall we see sword nor famine:

dixerunt non est ipse neque ueniet super nos malum gladium et famem non uidebimus 5:13 prophetae fuerunt in uentum 5:13 And the prophets shall become wind, and the word is

not in them: thus shall it be done unto them.

et responsum non fuit in eis haec ergo euenient illis 5:14 5:14 Wherefore thus saith the LORD God of hosts, Be-

cause ye speak this word, behold, I will make my words in thy mouth fire, and this people wood, and it shall devour

haec dicit Dominus Deus exercituum quia locuti estis uer- them. bum istud ecce ego do uerba mea in ore tuo in ignem et populum istum ligna et uorabit eos 5:15 ecce ego adducam 5:15 Lo, I will bring a nation upon you from far, O house

of Israel, saith the LORD: it is a mighty nation, it is an ancient nation, a nation whose language thou knowest not,

super uos gentem de longinquo domus Israhel ait Dominus neither understandest what they say. gentem robustam gentem antiquam gentem cuius ignorabis linguam nec intelleges quid loquatur 5:16 faretra eius quasi 5:16 Their quiver is as an open sepulchre, they are all

mighty men.

sepulchrum patens uniuersi fortes 5:17 et comedet segetes 5:17 And they shall eat up thine harvest, and thy bread,

which thy sons and thy daughters should eat: they shall eat up thy flocks and thine herds: they shall eat up thy vines and thy fig trees: they shall impoverish thy fenced cities, wherein thou trustedst, with the sword.

tuas et panem tuum deuorabit filios tuos et filias tuas comedet gregem tuum et armenta tua comedet uineam tuam et ficum tuam et conteret urbes munitas tuas in quibus tu habes fiduciam gladio 5:18 uerumtamen et diebus illis ait 5:18 Nevertheless in those days, saith the LORD, I will

not make a full end with you.

Dominus non faciam uos in consummationem 5:19 quod si 5:19 And it shall come to pass, when ye shall say, Wherefore doeth the LORD our God all these things unto us? then shalt thou answer them, Like as ye have forsaken me, and served strange gods in your land, so shall ye serve strangers in a land that is not your’s.

dixeritis quare fecit Dominus Deus noster nobis haec omnia dices ad eos sicut dereliquistis me et seruistis deo alieno in terra uestra sic seruietis alienis in terra non uestra 5:20 5:20 Declare this in the house of Jacob, and publish it in

Judah, saying,

adnuntiate hoc domui Iacob et auditum facite in Iuda dicentes 5:21 audi populus stulte qui non habes cor qui habentes 5:21 Hear now this, O foolish people, and without under-

standing; which have eyes, and see not; which have ears, and hear not:

oculos non uidetis et aures et non auditis 5:22 me ergo non

5:22 Fear ye not me? saith the LORD: will ye not tremble at my presence, which have placed the sand for the bound of the sea by a perpetual decree, that it cannot pass it: and though the waves thereof toss themselves, yet can they not prevail; though they roar, yet can they not pass over it?

Biblia Sacra Vulgata

timebitis ait Dominus et a facie mea non dolebitis qui posui harenam terminum mari praeceptum sempiternum quod non praeteribit et commouebuntur et non poterunt et intu-

5:23 But this people hath a revolting and a rebellious heart; mescent fluctus eius et non transibunt illud 5:23 populo authey are revolted and gone.

tem huic factum est cor incredulum et exasperans recesse-

5:24 Neither say they in their heart, Let us now fear the runt et abierunt 5:24 et non dixerunt in corde suo metua-

LORD our God, that giveth rain, both the former and the latter, in his season: he reserveth unto us the appointed

weeks of the harvest. mus Dominum Deum nostrum qui dat nobis pluuiam tem-

poraneam et serotinam in tempore suo plenitudinem annuae

5:25 Your iniquities have turned away these things, and messis custodientem nobis 5:25 iniquitates nostrae decli-

your sins have withholden good things from you.

nauerunt haec et peccata nostra prohibuerunt bonum a no-

5:26 For among my people are found wicked men: they bis 5:26 quia inuenti sunt in populo meo impii insidiantes

lay wait, as he that setteth snares; they set a trap, they catch men.

quasi aucupes laqueos ponentes et pedicas ad capiendos ui-

5:27 As a cage is full of birds, so are their houses full of ros 5:27 sicut decipula plena auibus sic domus eorum ple-

deceit: therefore they are become great, and waxen rich.

5:28 They are waxen fat, they shine: yea, they overpass the nae dolo ideo magnificati sunt et ditati 5:28 incrassati sunt

deeds of the wicked: they judge not the cause, the cause of the fatherless, yet they prosper; and the right of the needy

do they not judge. et inpinguati et praeterierunt sermones meos pessime cau-

sam non iudicauerunt causam pupilli non direxerunt et iudi-

5:29 Shall I not visit for these things? saith the LORD: cium pauperum non iudicauerunt 5:29 numquid super his shall not my soul be avenged on such a nation as this?

non uisitabo dicit Dominus aut super gentem huiuscemodi

5:30 A wonderful and horrible thing is committed in the non ulciscetur anima mea 5:30 stupor et mirabilia facta sunt

land;

5:31 The prophets prophesy falsely, and the priests bear in terra 5:31 prophetae prophetabant mendacium et sacer-

rule by their means; and my people love to have it so: and what will ye do in the end thereof?

dotes adplaudebant manibus suis et populus meus dilexit talia quid igitur fiet in nouissimo eius

6:1 O ye children of Benjamin, gather yourselves to flee 6:1 confortamini filii Beniamin in medio Hierusalem et in

out of the midst of Jerusalem, and blow the trumpet in Tekoa, and set up a sign of fire in Bethhaccerem: for evil

appeareth out of the north, and great destruction. Thecua clangite bucina et super Bethaccharem leuate uexil-

lum quia malum uisum est ab aquilone et contritio magna

6:2 I have likened the daughter of Zion to a comely and 6:2 speciosae et delicatae adsimilaui filiam Sion 6:3 ad eam

delicate woman. 6:3 The shepherds with their flocks shall come unto her; they shall pitch their tents against her round about; they shall feed every one in his place.

uenient pastores et greges eorum fixerunt in ea tentoria in

6:4 Prepare ye war against her; arise, and let us go up circuitu pascet unusquisque eos qui sub manu sua sunt 6:4

at noon. Woe unto us! for the day goeth away, for the shadows of the evening are stretched out.

sanctificate super eam bellum consurgite et ascendamus in meridie uae nobis quia declinauit dies quia longiores factae

6:5 Arise, and let us go by night, and let us destroy her sunt umbrae uesperi 6:5 surgite et ascendamus in nocte et

palaces.

6:6 For thus hath the LORD of hosts said, Hew ye down dissipemus domos eius 6:6 quia haec dicit Dominus exer-

trees, and cast a mount against Jerusalem: this is the city to be visited; she is wholly oppression in the midst of her.

cituum caedite lignum eius et fundite circa Hierusalem ag-

Prophetia Ieremiae

gerem haec est ciuitas uisitationis omnis calumnia in medio eius 6:7 sicut frigidam facit cisterna aquam suam sic fri- 6:7 As a fountain casteth out her waters, so she casteth out

her wickedness: violence and spoil is heard in her; before me continually is grief and wounds.

gidam fecit malitiam suam iniquitas et uastitas audietur in

ea coram me semper infirmitas et plaga 6:8 erudire Hie- 6:8 Be thou instructed, O Jerusalem, lest my soul depart

from thee; lest I make thee desolate, a land not inhabited.

rusalem ne forte recedat anima mea a te ne forte ponam te desertam terram inhabitabilem 6:9 haec dicit Dominus ex- 6:9 Thus saith the LORD of hosts, They shall throughly

glean the remnant of Israel as a vine: turn back thine hand as a grapegatherer into the baskets.

ercituum usque ad racemum colligent quasi in uinea reliquias Israhel conuerte manum tuam quasi uindemiator ad cartallum 6:10 cui loquar et quem contestabor ut audiant 6:10 To whom shall I speak, and give warning, that they

may hear? behold, their ear is uncircumcised, and they cannot hearken: behold, the word of the LORD is unto

ecce incircumcisae aures eorum et audire non possunt ecce them a reproach; they have no delight in it. uerbum Domini factum est eis in obprobrium et non suscipient illud 6:11 idcirco furore Domini plenus sum labo- 6:11 Therefore I am full of the fury of the LORD; I am

weary with holding in: I will pour it out upon the children abroad, and upon the assembly of young men together: for even the husband with the wife shall be taken, the aged with him that is full of days.

raui sustinens effunde super paruulum foris et super concilium iuuenum simul uir enim cum muliere capietur senex cum pleno dierum 6:12 et transibunt domus eorum ad alte- 6:12 And their houses shall be turned unto others, with

their fields and wives together: for I will stretch out my hand upon the inhabitants of the land, saith the LORD.

ros agri et uxores pariter quia extendam manum meam super habitantes terram dicit Dominus 6:13 a minore quippe 6:13 For from the least of them even unto the greatest of

them every one is given to covetousness; and from the prophet even unto the priest every one dealeth falsely.

usque ad maiorem omnes auaritiae student et a propheta usque ad sacerdotem cuncti faciunt dolum 6:14 et curabant 6:14 They have healed also the hurt of the daughter of

my people slightly, saying, Peace, peace; when there is no peace.

contritionem filiae populi mei cum ignominia dicentes pax pax et non erat pax 6:15 confusi sunt quia abominationem 6:15 Were they ashamed when they had committed abom-

ination? nay, they were not at all ashamed, neither could they blush: therefore they shall fall among them that fall: at the time that I visit them they shall be cast down, saith the LORD.

fecerunt quin potius confusione non sunt confusi et erubescere nescierunt quam ob rem cadent inter ruentes in tempore uisitationis suae corruent dicit Dominus 6:16 haec di- 6:16 Thus saith the LORD, Stand ye in the ways, and see,

and ask for the old paths, where is the good way, and walk therein, and ye shall find rest for your souls. But they said,

cit Dominus state super uias et uidete et interrogate de se- We will not walk therein. mitis antiquis quae sit uia bona et ambulate in ea et inuenietis refrigerium animabus uestris et dixerunt non ambulabimus 6:17 et constitui super uos speculatores audite uo- 6:17 Also I set watchmen over you, saying, Hearken to the

sound of the trumpet. But they said, We will not hearken.

cem tubae et dixerunt non audiemus 6:18 ideo audite gen- 6:18 Therefore hear, ye nations, and know, O congrega-

tion, what is among them.

tes et cognosce congregatio quanta ego faciam eis 6:19 audi 6:19 Hear, O earth: behold, I will bring evil upon this peo-

ple, even the fruit of their thoughts, because they have not hearkened unto my words, nor to my law, but rejected it.

terra ecce ego adducam mala super populum istum fructum cogitationum eius quia uerba mea non audierunt et legem meam proiecerunt 6:20 ut quid mihi tus de Saba adfertis 6:20 To what purpose cometh there to me incense from

Sheba, and the sweet cane from a far country? your burnt offerings are not acceptable, nor your sacrifices sweet unto me.

Biblia Sacra Vulgata

et calamum suaue olentem de terra longinqua holocaustomata uestra non sunt accepta et uictimae uestrae non pla-

6:21 Therefore thus saith the LORD, Behold, I will lay cuerunt mihi 6:21 propterea haec dicit Dominus ecce ego

stumblingblocks before this people, and the fathers and the sons together shall fall upon them; the neighbour and

his friend shall perish. dabo in populum istum ruinas et ruent in eis patres et filii 6:22 Thus saith the LORD, Behold, a people cometh from simul uicinus et proximus et peribunt 6:22 haec dicit Do-

the north country, and a great nation shall be raised from the sides of the earth.

minus ecce populus uenit de terra aquilonis et gens magna

6:23 They shall lay hold on bow and spear; they are cruel, consurget a finibus terrae 6:23 sagittam et scutum arripiet

and have no mercy; their voice roareth like the sea; and they ride upon horses, set in array as men for war against

thee, O daughter of Zion. crudelis est et non miserebitur uox eius quasi mare sonabit

et super equos ascendent praeparati quasi uir ad proelium ad-

6:24 We have heard the fame thereof: our hands wax fee- uersum te filia Sion 6:24 audiuimus famam eius dissolutae

ble: anguish hath taken hold of us, and pain, as of a woman in travail.

sunt manus nostrae tribulatio adprehendit nos dolores ut par-

6:25 Go not forth into the field, nor walk by the way; for turientem 6:25 nolite exire ad agros et in uia ne ambuletis

the sword of the enemy and fear is on every side.

6:26 O daughter of my people, gird thee with sackcloth, quoniam gladius inimici pauor in circuitu 6:26 filia populi

and wallow thyself in ashes: make thee mourning, as for an only son, most bitter lamentation: for the spoiler shall

suddenly come upon us. mei accingere cilicio et conspergere cinere luctum unigeniti

fac tibi planctum amarum quia repente ueniet uastator su-

6:27 I have set thee for a tower and a fortress among my per nos 6:27 probatorem dedi te in populo meo robustum

people, that thou mayest know and try their way. 6:28 They are all grievous revolters, walking with slan- et scies et probabis uiam eorum 6:28 omnes isti principes ders: they are brass and iron; they are all corrupters.

declinantum ambulantes fraudulenter aes et ferrum uniuersi

6:29 The bellows are burned, the lead is consumed of the corrupti sunt 6:29 defecit sufflatorium in igne consumptum

fire; the founder melteth in vain: for the wicked are not plucked away.

est plumbum frustra conflauit conflator malitiae enim eorum

6:30 Reprobate silver shall men call them, because the non sunt consumptae 6:30 argentum reprobum uocate eos LORD hath rejected them.

quia Dominus proiecit illos

7:1 The word that came to Jeremiah from the LORD, say- 7:1 uerbum quod factum est ad Hieremiam a Domino diing,

7:2 Stand in the gate of the LORD’s house, and proclaim cens 7:2 sta in porta domus Domini et praedica ibi uerbum

there this word, and say, Hear the word of the LORD, all ye of Judah, that enter in at these gates to worship the

LORD. istud et dic audite uerbum Domini omnis Iuda qui ingredi-7:3 Thus saith the LORD of hosts, the God of Israel, mini per portas has ut adoretis Dominum 7:3 haec dicit Do-

Amend your ways and your doings, and I will cause you to dwell in this place.

minus exercituum Deus Israhel bonas facite uias uestras et

7:4 Trust ye not in lying words, saying, The temple of studia uestra et habitabo uobiscum in loco isto 7:4 nolite

the LORD, The temple of the LORD, The temple of the LORD, are these.

confidere in uerbis mendacii dicentes templum Domini tem-

7:5 For if ye throughly amend your ways and your doings; plum Domini templum Domini est 7:5 quoniam si bene diif ye throughly execute judgment between a man and his neighbour;

rexeritis uias uestras et studia uestra si feceritis iudicium in-

7:6 If ye oppress not the stranger, the fatherless, and the ter uirum et proximum eius 7:6 aduenae et pupillo et uiduae widow, and shed not innocent blood in this place, neither walk after other gods to your hurt:

non feceritis calumniam nec sanguinem innocentem effude-

Prophetia Ieremiae

ritis in loco hoc et post deos alienos non ambulaueritis in

malum uobismet ipsis 7:7 habitabo uobiscum in loco isto 7:7 Then will I cause you to dwell in this place, in the land

that I gave to your fathers, for ever and ever.

in terra quam dedi patribus uestris a saeculo usque in saeculum 7:8 ecce uos confiditis uobis in sermonibus mendacii 7:8 Behold, ye trust in lying words, that cannot profit. qui non prodeeunt uobis 7:9 furari occidere adulterare iu- 7:9 Will ye steal, murder, and commit adultery, and swear

falsely, and burn incense unto Baal, and walk after other gods whom ye know not;

rare mendaciter libare Baali et ire post deos alienos quos ignoratis 7:10 et uenistis et stetistis coram me in domo hac in 7:10 And come and stand before me in this house, which

is called by my name, and say, We are delivered to do all these abominations?

qua inuocatum est nomen meum et dixistis liberati sumus eo quod fecerimus omnes abominationes istas 7:11 ergo spe- 7:11 Is this house, which is called by my name, become a

den of robbers in your eyes? Behold, even I have seen it, saith the LORD.

lunca latronum facta est domus ista in qua inuocatum est nomen meum in oculis uestris ego ego sum ego uidi dicit Dominus 7:12 ite ad locum meum in Silo ubi habitauit no- 7:12 But go ye now unto my place which was in Shiloh,

where I set my name at the first, and see what I did to it for the wickedness of my people Israel.

men meum a principio et uidete quae fecerim ei propter malitiam populi mei Israhel 7:13 et nunc quia fecistis omnia 7:13 And now, because ye have done all these works,

saith the LORD, and I spake unto you, rising up early and speaking, but ye heard not; and I called you, but ye an-

opera haec dicit Dominus et locutus sum ad uos mane con- swered not; surgens et loquens et non audistis et uocaui uos et non respondistis 7:14 faciam domui huic in qua inuocatum est no- 7:14 Therefore will I do unto this house, which is called

by my name, wherein ye trust, and unto the place which I gave to you and to your fathers, as I have done to Shiloh.

men meum et in qua uos habetis fiduciam et loco quem dedi uobis et patribus uestris sicut feci Silo 7:15 et proiciam 7:15 And I will cast you out of my sight, as I have cast out

all your brethren, even the whole seed of Ephraim.

uos a facie mea sicut proieci omnes fratres uestros uniuer-

sum semen Ephraim 7:16 tu ergo noli orare pro populo hoc 7:16 Therefore pray not thou for this people, neither lift up cry nor prayer for them, neither make intercession to me: for I will not hear thee.

nec adsumas pro eis laudem et orationem et non obsistas

mihi quia non exaudiam te 7:17 nonne uides quid isti fa- 7:17 Seest thou not what they do in the cities of Judah and

in the streets of Jerusalem?

ciant in ciuitatibus Iuda et in plateis Hierusalem 7:18 filii 7:18 The children gather wood, and the fathers kindle the

fire, and the women knead their dough, to make cakes to the queen of heaven, and to pour out drink offerings unto

colligunt ligna et patres succendunt ignem et mulieres con- other gods, that they may provoke me to anger. spergunt adipem ut faciant placentas Reginae caeli et libent diis alienis et me ad iracundiam prouocent 7:19 numquid 7:19 Do they provoke me to anger? saith the LORD: do

they not provoke themselves to the confusion of their own faces?

me ad iracundiam prouocant dicit Dominus nonne semet ipsos in confusionem uultus sui 7:20 ideo haec dicit Domi- 7:20 Therefore thus saith the Lord GOD; Behold, mine

anger and my fury shall be poured out upon this place, upon man, and upon beast, and upon the trees of the field, and upon the fruit of the ground; and it shall burn, and shall not be quenched.

nus Deus ecce furor meus et indignatio mea conflatur super locum istum super uiros et super iumenta et super lignum regionis et super fruges terrae et succendetur et non extinguetur 7:21 haec dicit Dominus exercituum Deus Israhel ho- 7:21 Thus saith the LORD of hosts, the God of Israel; Put

your burnt offerings unto your sacrifices, and eat flesh.

Biblia Sacra Vulgata

locaustomata uestra addite uictimis uestris et comedite car-

7:22 For I spake not unto your fathers, nor commanded nes 7:22 quia non sum locutus cum patribus uestris et non

them in the day that I brought them out of the land of Egypt, concerning burnt offerings or sacrifices:

praecepi eis in die qua eduxi eos de terra Aegypti de uerbo

7:23 But this thing commanded I them, saying, Obey my holocaustomatum et uictimarum 7:23 sed hoc uerbum prae-

voice, and I will be your God, and ye shall be my people: and walk ye in all the ways that I have commanded you,

that it may be well unto you. cepi eis dicens audite uocem meam et ero uobis Deus et uos

eritis mihi populus et ambulate in omni uia quam mandaui

7:24 But they hearkened not, nor inclined their ear, but uobis ut bene sit uobis 7:24 et non audierunt nec inclina-

walked in the counsels and in the imagination of their evil heart, and went backward, and not forward.

uerunt aurem suam sed abierunt in uoluntatibus et prauitate

7:25 Since the day that your fathers came forth out of the cordis sui mali factique sunt retrorsum et non in ante 7:25

land of Egypt unto this day I have even sent unto you all my servants the prophets, daily rising up early and sending

them: a die qua egressi sunt patres eorum de terra Aegypti usque

ad diem hanc et misi ad uos omnes seruos meos prophe-

7:26 Yet they hearkened not unto me, nor inclined their tas per diem consurgens diluculo et mittens 7:26 et non au-

ear, but hardened their neck: they did worse than their fathers.

dierunt me nec inclinauerunt aurem suam sed indurauerunt

7:27 Therefore thou shalt speak all these words unto them; ceruicem et peius operati sunt quam patres eorum 7:27 et but they will not hearken to thee: thou shalt also call unto them; but they will not answer thee.

loqueris ad eos omnia uerba haec et non audient te et uoca-

7:28 But thou shalt say unto them, This is a nation that bis eos et non respondebunt tibi 7:28 et dices ad eos haec

obeyeth not the voice of the LORD their God, nor receiveth correction: truth is perished, and is cut off from

their mouth. est gens quae non audiuit uocem Domini Dei sui nec rece-7:29 Cut off thine hair, O Jerusalem, and cast it away, and pit disciplinam periit fides et ablata est de ore eorum 7:29

take up a lamentation on high places; for the LORD hath rejected and forsaken the generation of his wrath.

tonde capillum tuum et proice et sume in directum planctum quia proiecit Dominus et reliquit generationem furoris sui

7:30 For the children of Judah have done evil in my sight, 7:30 quia fecerunt filii Iuda malum in oculis meis dicit Do-

saith the LORD: they have set their abominations in the house which is called by my name, to pollute it.

minus posuerunt offendicula sua in domo in qua inuocatum

7:31 And they have built the high places of Tophet, which est nomen meum ut polluerent eam 7:31 et aedificauerunt is in the valley of the son of Hinnom, to burn their sons and their daughters in the fire; which I commanded them

not, neither came it into my heart. excelsa Thofeth qui est in ualle filii Ennom ut incenderent fi-

lios suos et filias suas igni quae non praecepi nec cogitaui in

7:32 Therefore, behold, the days come, saith the LORD, corde meo 7:32 ideo ecce dies uenient dicit Dominus et non

that it shall no more be called Tophet, nor the valley of the son of Hinnom, but the valley of slaughter: for they shall

bury in Tophet, till there be no place. dicetur amplius Thofeth et uallis filii Ennom sed uallis Inter-7:33 And the carcases of this people shall be meat for the fectionis et sepelient in Thofeth eo quod non sit locus 7:33

fowls of the heaven, and for the beasts of the earth; and none shall fray them away.

et erit morticinum populi huius in cibum uolucribus caeli et

7:34 Then will I cause to cease from the cities of Judah, bestiis terrae et non erit qui abigat 7:34 et quiescere faciam

and from the streets of Jerusalem, the voice of mirth, and the voice of gladness, the voice of the bridegroom, and the

voice of the bride: for the land shall be desolate. de urbibus Iuda et de plateis Hierusalem uocem gaudii et uocem laetitiae uocem sponsi et uocem sponsae in desolatione enim erit terra

Prophetia Ieremiae

8:1 in tempore illo ait Dominus eicient ossa regis Iuda et 8:1 At that time, saith the LORD, they shall bring out the

bones of the kings of Judah, and the bones of his princes, and the bones of the priests, and the bones of the prophets, and the bones of the inhabitants of Jerusalem, out of their graves:

ossa principum eius et ossa sacerdotum et ossa prophetarum et ossa eorum qui habitauerunt Hierusalem de sepulchris suis 8:2 et pandent ea ad solem et lunam et omnem 8:2 And they shall spread them before the sun, and the

moon, and all the host of heaven, whom they have loved, and whom they have served, and after whom they have walked, and whom they have sought, and whom they have worshipped: they shall not be gathered, nor be buried; they shall be for dung upon the face of the earth.

militiam caeli quae dilexerunt et quibus seruierunt et post quae ambulauerunt et quae quaesierunt et adorauerunt non colligentur et non sepelientur in sterquilinium super faciem terrae erunt 8:3 et eligent magis mortem quam uitam om- 8:3 And death shall be chosen rather than life by all the

residue of them that remain of this evil family, which remain in all the places whither I have driven them, saith the

nes qui residui fuerint de cognatione hac pessima in uniuer- LORD of hosts. sis locis quae derelicta sunt ad quae eieci eos dicit Dominus exercituum 8:4 et dices ad eos haec dicit Dominus num- 8:4 Moreover thou shalt say unto them, Thus saith the

LORD; Shall they fall, and not arise? shall he turn away, and not return?

quid qui cadet non resurget et qui auersus est non reuertetur 8:5 quare ergo auersus est populus iste in Hierusalem 8:5 Why then is this people of Jerusalem slidden back by

a perpetual backsliding? they hold fast deceit, they refuse to return.

auersione contentiosa adprehenderunt mendacium et noluerunt reuerti 8:6 adtendi et auscultaui nemo quod bonum est 8:6 I hearkened and heard, but they spake not aright: no

man repented him of his wickedness, saying, What have I done? every one turned to his course, as the horse rusheth

loquitur nullus est qui agat paenitentiam super peccato suo into the battle. dicens quid feci omnes conuersi sunt ad cursum suum quasi equus impetu uadens in proelio 8:7 miluus in caelo cog- 8:7 Yea, the stork in the heaven knoweth her appointed

times; and the turtle and the crane and the swallow observe the time of their coming; but my people know not

nouit tempus suum turtur et hirundo et ciconia custodierunt the judgment of the LORD. tempus aduentus sui populus autem meus non cognouit iudicium Domini 8:8 quomodo dicitis sapientes nos sumus et 8:8 How do ye say, We are wise, and the law of the LORD

is with us? Lo, certainly in vain made he it; the pen of the scribes is in vain.

lex Domini nobiscum est uere mendacium operatus est stilus mendax scribarum 8:9 confusi sunt sapientes perterriti 8:9 The wise men are ashamed, they are dismayed and

taken: lo, they have rejected the word of the LORD; and what wisdom is in them?

et capti sunt uerbum enim Domini proiecerunt et sapientia

nulla est in eis 8:10 propterea dabo mulieres eorum exteris 8:10 Therefore will I give their wives unto others, and their fields to them that shall inherit them: for every one from the least even unto the greatest is given to covetousness, from the prophet even unto the priest every one dealeth falsely.

agros eorum heredibus quia a minimo usque ad maximum omnes auaritiam sequuntur a propheta usque ad sacerdotem cuncti faciunt mendacium 8:11 et sanabant contritionem fi- 8:11 For they have healed the hurt of the daughter of my

people slightly, saying, Peace, peace; when there is no peace.

liae populi mei ad ignominiam dicentes pax pax cum non esset pax 8:12 confusi sunt quia abominationem fecerunt 8:12 Were they ashamed when they had committed abom-

ination? nay, they were not at all ashamed, neither could they blush: therefore shall they fall among them that fall: in the time of their visitation they shall be cast down, saith the LORD.

quinimmo confusione non sunt confusi et erubescere nescierunt idcirco cadent inter corruentes in tempore uisitationis suae corruent dicit Dominus 8:13 congregans congre- 8:13 I will surely consume them, saith the LORD: there

shall be no grapes on the vine, nor figs on the fig tree, and the leaf shall fade; and the things that I have given them shall pass away from them.

Biblia Sacra Vulgata

gabo eos ait Dominus non est uua in uitibus et non sunt ficus in ficulnea folium defluxit et dedi eis quae praetergressa

8:14 Why do we sit still? assemble yourselves, and let us sunt 8:14 quare sedemus conuenite et ingrediamur ciuita-

enter into the defenced cities, and let us be silent there: for the LORD our God hath put us to silence, and given us water of gall to drink, because we have sinned against the LORD.

tem munitam et sileamus ibi quia Dominus noster silere nos fecit et potum dedit nobis aquam fellis peccauimus enim Do-

8:15 We looked for peace, but no good came; and for a mino 8:15 expectauimus pacem et non erat bonum tempus

time of health, and behold trouble!

8:16 The snorting of his horses was heard from Dan: the medellae et ecce formido 8:16 a Dan auditus est fremitus

whole land trembled at the sound of the neighing of his strong ones; for they are come, and have devoured the land, and all that is in it; the city, and those that dwell therein.

equorum eius a uoce hinnituum pugnatorum eius commota est omnis terra et uenerunt et deuorauerunt terram et pleni-

8:17 For, behold, I will send serpents, cockatrices, among tudinem eius urbem et habitatores eius 8:17 quia ecce ego you, which will not be charmed, and they shall bite you, saith the LORD.

mittam uobis serpentes regulos quibus non est incantatio et

8:18 When I would comfort myself against sorrow, my mordebunt uos ait Dominus 8:18 dolor meus super dolo-

heart is faint in me.

8:19 Behold the voice of the cry of the daughter of my rem in me cor meum maerens 8:19 ecce uox clamoris filiae

people because of them that dwell in a far country: Is not the LORD in Zion? is not her king in her? Why have they provoked me to anger with their graven images, and with strange vanities?

populi mei de terra longinqua numquid Dominus non est in Sion aut rex eius non est in ea quare ergo me ad iracundiam concitauerunt in sculptilibus suis et in uanitatibus alie-

8:20 The harvest is past, the summer is ended, and we are nis 8:20 transiit messis finita est aestas et nos saluati non

not saved.

8:21 For the hurt of the daughter of my people am I hurt; sumus 8:21 super contritionem filiae populi mei contritus

I am black; astonishment hath taken hold on me.

8:22 Is there no balm in Gilead; is there no physician sum et contristatus stupor obtinuit me 8:22 numquid resina

there? why then is not the health of the daughter of my people recovered?

non est in Galaad aut medicus non est ibi quare igitur non est obducta cicatrix filiae populi mei

9:1 Oh that my head were waters, and mine eyes a fountain 9:1 quis dabit capiti meo aquam et oculis meis fontem lac-

of tears, that I might weep day and night for the slain of the daughter of my people!

rimarum et plorabo die et nocte interfectos filiae populi mei

9:2 Oh that I had in the wilderness a lodging place of way- 9:2 quis dabit me in solitudine diuersorium uiatorum et de-

faring men; that I might leave my people, and go from them! for they be all adulterers, an assembly of treacher-

ous men. relinquam populum meum et recedam ab eis quia omnes ad-9:3 And they bend their tongues like their bow for lies: but ulteri sunt coetus praeuaricatorum 9:3 et extenderunt lin-

they are not valiant for the truth upon the earth; for they proceed from evil to evil, and they know not me, saith the

LORD. guam suam quasi arcum mendacii et non ueritatis confortati

sunt in terra quia de malo ad malum egressi sunt et me non

9:4 Take ye heed every one of his neighbour, and trust ye cognouerunt dicit Dominus 9:4 unusquisque se a proximo

not in any brother: for every brother will utterly supplant, and every neighbour will walk with slanders.

suo custodiat et in omni fratre suo non habeat fiduciam quia omnis frater subplantans subplantabit et omnis amicus frau-

9:5 And they will deceive every one his neighbour, and dulenter incedet 9:5 et uir fratrem suum deridebit et uerita-

will not speak the truth: they have taught their tongue to speak lies, and weary themselves to commit iniquity.

tem non loquentur docuerunt enim linguam suam loqui men-

Prophetia Ieremiae

dacium ut inique agerent laborauerunt 9:6 habitatio tua in 9:6 Thine habitation is in the midst of deceit; through de-

ceit they refuse to know me, saith the LORD.

medio doli in dolo rennuerunt scire me dicit Dominus 9:7 9:7 Therefore thus saith the LORD of hosts, Behold, I will melt them, and try them; for how shall I do for the daughter of my people?

propterea haec dicit Dominus exercituum ecce ego conflabo et probabo eos quid enim aliud faciam a facie filiae populi

mei 9:8 sagitta uulnerans lingua eorum dolum locuta est in 9:8 Their tongue is as an arrow shot out; it speaketh deceit: one speaketh peaceably to his neighbour with his mouth, but in heart he layeth his wait.

ore suo pacem cum amico suo loquitur et occulte ponit ei insidias 9:9 numquid super his non uisitabo dicit Dominus 9:9 Shall I not visit them for these things? saith the LORD:

shall not my soul be avenged on such a nation as this?

aut in gentem huiuscemodi non ulciscetur anima mea 9:10 9:10 For the mountains will I take up a weeping and wail-

ing, and for the habitations of the wilderness a lamentation, because they are burned up, so that none can pass through them; neither can men hear the voice of the cattle; both the fowl of the heavens and the beast are fled; they are gone.

super montes adsumam fletum ac lamentum et super speciosa deserti planctum quoniam incensa sunt eo quod non sit uir pertransiens et non audierunt uocem possidentis a uolucre caeli usque ad pecora transmigrauerunt et recesserunt 9:11 et dabo Hierusalem in aceruos harenae et cubilia dra- 9:11 And I will make Jerusalem heaps, and a den of drag-

ons; and I will make the cities of Judah desolate, without an inhabitant.

conum et ciuitates Iuda dabo in desolationem eo quod non sit habitator 9:12 quis est uir sapiens qui intellegat hoc et ad 9:12 Who is the wise man, that may understand this? and

who is he to whom the mouth of the LORD hath spoken, that he may declare it, for what the land perisheth and is

quem uerbum oris Domini fiat ut adnuntiet istud quare perie- burned up like a wilderness, that none passeth through? rit terra exusta sit quasi desertum eo quod non sit qui pertranseat 9:13 et dixit Dominus quia dereliquerunt legem meam 9:13 And the LORD saith, Because they have forsaken

my law which I set before them, and have not obeyed my voice, neither walked therein;

quam dedi eis et non audierunt uocem meam et non ambulauerunt in ea 9:14 et abierunt post prauitatem cordis sui 9:14 But have walked after the imagination of their own

heart, and after Baalim, which their fathers taught them:

et post Baalim quos didicerunt a patribus suis 9:15 idcirco 9:15 Therefore thus saith the LORD of hosts, the God of

Israel; Behold, I will feed them, even this people, with wormwood, and give them water of gall to drink.

haec dicit Dominus exercituum Deus Israhel ecce ego cibabo eos populum istum absinthio et potum dabo eis aquam fellis 9:16 et dispergam eos in gentibus quas non nouerunt ipsi et 9:16 I will scatter them also among the heathen, whom

neither they nor their fathers have known: and I will send a sword after them, till I have consumed them.

patres eorum et mittam post eos gladium donec consumantur 9:17 haec dicit Dominus exercituum contemplamini et uocate 9:17 Thus saith the LORD of hosts, Consider ye, and call

for the mourning women, that they may come; and send for cunning women, that they may come:

lamentatrices et ueniant et ad eas quae sapientes sunt mittite et properent 9:18 festinent et adsumant super nos lamentum 9:18 And let them make haste, and take up a wailing for

us, that our eyes may run down with tears, and our eyelids gush out with waters.

deducant oculi nostri lacrimas et palpebrae nostrae defluant aquis 9:19 quia uox lamentationis audita est de Sion quo- 9:19 For a voice of wailing is heard out of Zion, How are

we spoiled! we are greatly confounded, because we have forsaken the land, because our dwellings have cast us out.

modo uastati sumus et confusi uehementer quia dereliquimus terram quoniam deiecta sunt tabernacula nostra 9:20 9:20 Yet hear the word of the LORD, O ye women, and

let your ear receive the word of his mouth, and teach your daughters wailing, and every one her neighbour lamenta-

audite ergo mulieres uerbum Domini et adsumat auris uestra tion.

Biblia Sacra Vulgata

sermonem oris eius et docete filias uestras lamentum et una-

9:21 For death is come up into our windows, and is entered quaeque proximam suam planctum 9:21 quia ascendit mors into our palaces, to cut off the children from without, and the young men from the streets.

per fenestras nostras ingressa est domos nostras disperdere

9:22 Speak, Thus saith the LORD, Even the carcases of paruulos de foris iuuenes de plateis 9:22 loquere haec di-

men shall fall as dung upon the open field, and as the handful after the harvestman, and none shall gather them.

cit Dominus et cadet morticinum hominis quasi stercus super faciem regionis et quasi faenum post tergum metentis et

9:23 Thus saith the LORD, Let not the wise man glory in non est qui colligat 9:23 haec dicit Dominus non glorietur

his wisdom, neither let the mighty man glory in his might, let not the rich man glory in his riches:

sapiens in sapientia sua et non glorietur fortis in fortitudine

9:24 But let him that glorieth glory in this, that he un- sua et non glorietur diues in diuitiis suis 9:24 sed in hoc glo-

derstandeth and knoweth me, that I am the LORD which exercise lovingkindness, judgment, and righteousness, in

the earth: for in these things I delight, saith the LORD. rietur qui gloriatur scire et nosse me quia ego sum Dominus

qui facio misericordiam et iudicium et iustitiam in terra haec

9:25 Behold, the days come, saith the LORD, that I will enim placent mihi ait Dominus 9:25 ecce dies ueniunt dicit

punish all them which are circumcised with the uncircumcised;

Dominus et uisitabo super omnem qui circumcisum habet

9:26 Egypt, and Judah, and Edom, and the children of Am- praeputium 9:26 super Aegyptum et super Iudam et super mon, and Moab, and all that are in the utmost corners, that dwell in the wilderness: for all these nations are uncircumcised, and all the house of Israel are uncircumcised in the heart.

Edom et super filios Ammon et super Moab et super omnes qui adtonsi sunt in comam habitantes in deserto quia omnes gentes habent praeputium omnis autem domus Israhel incircumcisi sunt corde

10:1 Hear ye the word which the LORD speaketh unto 10:1 audite uerbum quod locutus est Dominus super uos

you, O house of Israel:

10:2 Thus saith the LORD, Learn not the way of the hea- domus Israhel 10:2 haec dicit Dominus iuxta uias gentium

then, and be not dismayed at the signs of heaven; for the heathen are dismayed at them.

nolite discere et a signis caeli nolite metuere quae timent

10:3 For the customs of the people are vain: for one cut- gentes 10:3 quia leges populorum uanae sunt quia lignum

teth a tree out of the forest, the work of the hands of the workman, with the axe.

de saltu praecidit opus manuum artificis in ascia 10:4 ar-

10:4 They deck it with silver and with gold; they fasten it with nails and with hammers, that it move not.

gento et auro decorauit illud clauis et malleis conpegit ut non

10:5 They are upright as the palm tree, but speak not: they dissoluatur 10:5 in similitudinem palmae fabricata sunt et

must needs be borne, because they cannot go. Be not afraid of them; for they cannot do evil, neither also is it

in them to do good. non loquentur portata tollentur quia incedere non ualent no-

lite ergo timere ea quia nec male possunt facere nec bene

10:6 Forasmuch as there is none like unto thee, O LORD; 10:6 non est similis tui Domine magnus tu et magnum nothou art great, and thy name is great in might.

10:7 Who would not fear thee, O King of nations? for men tuum in fortitudine 10:7 quis non timebit te o rex gen-

to thee doth it appertain: forasmuch as among all the wise men of the nations, and in all their kingdoms, there is none

like unto thee. tium tuum est enim decus inter cunctos sapientes gentium et 10:8 But they are altogether brutish and foolish: the stock in uniuersis regnis eorum nullus est similis tui 10:8 pariter

is a doctrine of vanities.

insipientes et fatui probabuntur doctrina uanitatis eorum li-

10:9 Silver spread into plates is brought from Tarshish, gnum est 10:9 argentum inuolutum de Tharsis adfertur et and gold from Uphaz, the work of the workman, and of the hands of the founder: blue and purple is their clothing: they are all the work of cunning men.

Prophetia Ieremiae

aurum de Ofaz opus artificis et manus aerarii hyacinthus et purpura indumentum eorum opus artificum uniuersa haec 10:10 Dominus autem Deus uerus est ipse Deus uiuens et rex 10:10 But the LORD is the true God, he is the living God,

and an everlasting king: at his wrath the earth shall tremble, and the nations shall not be able to abide his indigna-

sempiternus ab indignatione eius commouebitur terra et non tion. sustinebunt gentes comminationem eius 10:11 sic ergo di- 10:11 Thus shall ye say unto them, The gods that have

not made the heavens and the earth, even they shall perish from the earth, and from under these heavens.

cetis eis dii qui caelos et terram non fecerunt pereant de terra et de his quae sub caelis sunt 10:12 qui facit terram in for- 10:12 He hath made the earth by his power, he hath estab-

lished the world by his wisdom, and hath stretched out the heavens by his discretion.

titudine sua praeparat orbem in sapientia sua et prudentia sua extendit caelos 10:13 ad uocem suam dat multitudinem 10:13 When he uttereth his voice, there is a multitude of

waters in the heavens, and he causeth the vapours to ascend from the ends of the earth; he maketh lightnings with

aquarum in caelo et eleuat nebulas ab extremitatibus terrae rain, and bringeth forth the wind out of his treasures. fulgura in pluuiam facit et educit uentum de thesauris suis 10:14 stultus factus est omnis homo ab scientia confusus est 10:14 Every man is brutish in his knowledge: every

founder is confounded by the graven image: for his molten image is falsehood, and there is no breath in them.

omnis artifex in sculptili quoniam falsum est quod conflauit et non est spiritus in eis 10:15 uana sunt et opus risu dig- 10:15 They are vanity, and the work of errors: in the time

of their visitation they shall perish.

num in tempore uisitationis suae peribunt 10:16 non est his 10:16 The portion of Jacob is not like them: for he is the

former of all things; and Israel is the rod of his inheritance: The LORD of hosts is his name.

similis pars Iacob qui enim formauit omnia ipse est et Israhel uirga hereditatis eius Dominus exercituum nomen illi 10:17 congrega de terra confusionem tuam quae habitas in 10:17 Gather up thy wares out of the land, O inhabitant of

the fortress.

obsidione 10:18 quia haec dicit Dominus ecce ego longe 10:18 For thus saith the LORD, Behold, I will sling out

the inhabitants of the land at this once, and will distress them, that they may find it so.

proiciam habitatores terrae in hac uice et tribulabo eos ita ut inueniantur 10:19 uae mihi super contritione mea pessima 10:19 Woe is me for my hurt! my wound is grievous; but

I said, Truly this is a grief, and I must bear it.

plaga mea ego autem dixi plane haec infirmitas mea est et portabo illam 10:20 tabernaculum meum uastatum est om- 10:20 My tabernacle is spoiled, and all my cords are bro-

ken: my children are gone forth of me, and they are not: there is none to stretch forth my tent any more, and to set

nes funiculi mei disrupti sunt filii mei exierunt a me et non up my curtains. subsistunt non est qui extendat ultra tentorium meum et erigat pelles meas 10:21 quia stulte egerunt pastores et Domi- 10:21 For the pastors are become brutish, and have not

sought the LORD: therefore they shall not prosper, and all their flocks shall be scattered.

num non quaesierunt propterea non intellexerunt et omnis grex eorum dispersus est 10:22 uox auditionis ecce uenit et 10:22 Behold, the noise of the bruit is come, and a great

commotion out of the north country, to make the cities of Judah desolate, and a den of dragons.

commotio magna de terra aquilonis ut ponat ciuitates Iuda solitudinem et habitaculum draconum 10:23 scio Domine 10:23 O LORD, I know that the way of man is not in him-

self: it is not in man that walketh to direct his steps.

quia non est hominis uia eius nec uiri est ut ambulet et dirigat gressus suos 10:24 corripe me Domine uerumtamen in 10:24 O LORD, correct me, but with judgment; not in

thine anger, lest thou bring me to nothing.

iudicio et non in furore tuo ne forte ad nihilum rediias me

Biblia Sacra Vulgata

10:25 Pour out thy fury upon the heathen that know thee 10:25 effunde indignationem tuam super gentes quae non co-

not, and upon the families that call not on thy name: for they have eaten up Jacob, and devoured him, and con-

sumed him, and have made his habitation desolate. gnouerunt te et super prouincias quae nomen tuum non inuocauerunt quia comederunt Iacob et deuorauerunt eum et consumpserunt illum et decus eius dissipauerunt

11:1 The word that came to Jeremiah from the LORD say- 11:1 uerbum quod factum est ad Hieremiam a Domino diing,

11:2 Hear ye the words of this covenant, and speak unto cens 11:2 audite uerba pacti huius et loquimini ad uiros

the men of Judah, and to the inhabitants of Jerusalem;

11:3 And say thou unto them, Thus saith the LORD God Iuda et habitatores Hierusalem 11:3 et dices ad eos haec

of Israel; Cursed be the man that obeyeth not the words of this covenant,

dicit Dominus Deus Israhel maledictus uir qui non audierit

11:4 Which I commanded your fathers in the day that I uerba pacti huius 11:4 quod praecepi patribus uestris in die

brought them forth out of the land of Egypt, from the iron furnace, saying, Obey my voice, and do them, according to all which I command you: so shall ye be my people, and I will be your God:

qua eduxi eos de terra Aegypti de fornace ferrea dicens audite uocem meam et facite omnia quae praecipio uobis et eri-

11:5 That I may perform the oath which I have sworn unto tis mihi in populum et ego ero uobis in Deum 11:5 ut susci-

your fathers, to give them a land flowing with milk and honey, as it is this day. Then answered I, and said, So be

it, O LORD. tem iuramentum quod iuraui patribus uestris daturum me eis

terram fluentem lacte et melle sicut est dies haec et respondi

11:6 Then the LORD said unto me, Proclaim all these et dixi amen Domine 11:6 et dixit Dominus ad me uocife-

words in the cities of Judah, and in the streets of Jerusalem, saying, Hear ye the words of this covenant, and

do them. rare omnia uerba haec in ciuitatibus Iuda et foris Hierusalem 11:7 For I earnestly protested unto your fathers in the day dicens audite uerba pacti huius et facite illa 11:7 quia con-

that I brought them up out of the land of Egypt, even unto this day, rising early and protesting, saying, Obey my

voice. testans contestatus sum patres uestros in die qua eduxi eos

de terra Aegypti usque ad diem hanc mane surgens contes-

11:8 Yet they obeyed not, nor inclined their ear, but walked tatus sum et dixi audite uocem meam 11:8 et non audierunt every one in the imagination of their evil heart: therefore I will bring upon them all the words of this covenant, which

I commanded them to do: but they did them not. nec inclinauerunt aurem suam sed abierunt unusquisque in

prauitate cordis sui mali et induxi super eos omnia uerba

11:9 And the LORD said unto me, A conspiracy is found pacti huius quod praecepi ut facerent et non fecerunt 11:9

among the men of Judah, and among the inhabitants of Jerusalem.

et dixit Dominus ad me inuenta est coniuratio in uiris Iuda

11:10 They are turned back to the iniquities of their fore- et in habitatoribus Hierusalem 11:10 reuersi sunt ad iniqui-

fathers, which refused to hear my words; and they went after other gods to serve them: the house of Israel and the house of Judah have broken my covenant which I made with their fathers.

tates patrum suorum priores qui noluerunt audire uerba mea et hii ergo abierunt post deos alienos ut seruirent eis irritum fecerunt domus Israhel et domus Iuda pactum meum quod

11:11 Therefore thus saith the LORD, Behold, I will bring pepigi cum patribus eorum 11:11 quam ob rem haec dicit evil upon them, which they shall not be able to escape; and though they shall cry unto me, I will not hearken unto

them. Dominus ecce ego inducam super eos mala de quibus exire 11:12 Then shall the cities of Judah and inhabitants of non poterunt et clamabunt ad me et non exaudiam eos 11:12

Jerusalem go, and cry unto the gods unto whom they offer incense: but they shall not save them at all in the time of

their trouble. et ibunt ciuitates Iuda et habitatores Hierusalem et clamabunt ad deos quibus libant et non saluabunt eos in tempore

Prophetia Ieremiae

adflictionis eorum 11:13 secundum numerum enim ciuita- 11:13 For according to the number of thy cities were thy

gods, O Judah; and according to the number of the streets of Jerusalem have ye set up altars to that shameful thing,

tum tuarum erant dii tui Iuda et secundum numerum uiarum even altars to burn incense unto Baal. Hierusalem posuistis aras confusionis aras ad libandum Ba-

ali 11:14 tu ergo noli orare pro populo hoc et ne adsumas 11:14 Therefore pray not thou for this people, neither lift up a cry or prayer for them: for I will not hear them in the time that they cry unto me for their trouble.

pro eis laudem et orationem quia non exaudiam in tempore clamoris eorum ad me in tempore adflictionis eorum 11:15 11:15 What hath my beloved to do in mine house, seeing

she hath wrought lewdness with many, and the holy flesh is passed from thee? when thou doest evil, then thou re-

quid est quod dilectus meus in domo mea fecit scelera multa joicest. numquid carnes sanctae auferent a te malitias tuas in quibus gloriata es 11:16 oliuam uberem pulchram fructiferam 11:16 The LORD called thy name, A green olive tree, fair,

and of goodly fruit: with the noise of a great tumult he hath kindled fire upon it, and the branches of it are broken.

speciosam uocauit Dominus nomen tuum ad uocem loquellae grandis exarsit ignis in ea et conbusta sunt frutecta eius 11:17 et Dominus exercituum qui plantauit te locutus est su- 11:17 For the LORD of hosts, that planted thee, hath pro-

nounced evil against thee, for the evil of the house of Israel and of the house of Judah, which they have done against themselves to provoke me to anger in offering incense unto Baal.

per te malum pro malis domus Israhel et domus Iuda quae fecerunt sibi ad inritandum me libantes Baali 11:18 tu au- 11:18 And the LORD hath given me knowledge of it, and

I know it: then thou shewedst me their doings.

tem Domine demonstrasti mihi et cognoui tunc ostendisti mihi studia eorum 11:19 et ego quasi agnus mansuetus qui 11:19 But I was like a lamb or an ox that is brought to the

slaughter; and I knew not that they had devised devices against me, saying, Let us destroy the tree with the fruit thereof, and let us cut him off from the land of the living, that his name may be no more remembered.

portatur ad uictimam et non cognoui quia super me cogitauerunt consilia mittamus lignum in panem eius et eradamus eum de terra uiuentium et nomen eius non memoretur am-

plius 11:20 tu autem Domine Sabaoth qui iudicas iuste et 11:20 But, O LORD of hosts, that judgest righteously, that triest the reins and the heart, let me see thy vengeance on them: for unto thee have I revealed my cause.

probas renes et cor uideam ultionem tuam ex eis tibi enim reuelaui causam meam 11:21 propterea haec dicit Domi- 11:21 Therefore thus saith the LORD of the men of

Anathoth, that seek thy life, saying, Prophesy not in the name of the LORD, that thou die not by our hand:

nus ad uiros Anathoth qui quaerunt animam tuam et dicunt non prophetabis in nomine Domini et non morieris in manibus nostris 11:22 propterea haec dicit Dominus exercituum 11:22 Therefore thus saith the LORD of hosts, Behold, I

will punish them: the young men shall die by the sword; their sons and their daughters shall die by famine:

ecce ego uisitabo super eos iuuenes morientur in gladio filii eorum et filiae eorum morientur in fame 11:23 et reliquiae 11:23 And there shall be no remnant of them: for I will

bring evil upon the men of Anathoth, even the year of their visitation.

non erunt ex eis inducam enim malum super uiros Anathoth annum uisitationis eorum

12:1 iustus quidem tu es Domine si disputem tecum ue- 12:1 Righteous art thou, O LORD, when I plead with thee: yet let me talk with thee of thy judgments: Wherefore doth the way of the wicked prosper? wherefore are all they

rumtamen iusta loquar ad te quare uia impiorum prosperatur happy that deal very treacherously? bene est omnibus qui praeuaricantur et inique agunt 12:2 12:2 Thou hast planted them, yea, they have taken root:

they grow, yea, they bring forth fruit: thou art near in their mouth, and far from their reins.

plantasti eos et radicem miserunt proficiunt et faciunt fruc-

Biblia Sacra Vulgata

12:3 But thou, O LORD, knowest me: thou hast seen me, tum prope es tu ori eorum et longe a renibus eorum 12:3

and tried mine heart toward thee: pull them out like sheep for the slaughter, and prepare them for the day of slaugh-

ter. et tu Domine nosti me uidisti me et probasti cor meum te-

cum congrega eos quasi gregem ad uictimam et sanctifica

12:4 How long shall the land mourn, and the herbs of every eos in die occisionis 12:4 usquequo lugebit terra et herba

field wither, for the wickedness of them that dwell therein? the beasts are consumed, and the birds; because they said,

He shall not see our last end. omnis regionis siccabitur propter malitiam habitantium in

ea consumptum est animal et uolucre quoniam dixerunt non

12:5 If thou hast run with the footmen, and they have wea- uidebit nouissima nostra 12:5 si cum peditibus currens la-

ried thee, then how canst thou contend with horses? and if in the land of peace, wherein thou trustedst, they wearied

thee, then how wilt thou do in the swelling of Jordan? borasti quomodo contendere poteris cum equis cum autem

in terra pacis secura fueris quid facies in superbia Iorda-

12:6 For even thy brethren, and the house of thy fa- nis 12:6 nam et fratres tui et domus patris tui etiam ipsi

ther, even they have dealt treacherously with thee; yea, they have called a multitude after thee: believe them not,

though they speak fair words unto thee. pugnauerunt aduersum te et clamauerunt post te plena uoce

12:7 I have forsaken mine house, I have left mine heritage; ne credas eis cum locuti fuerint tibi bona 12:7 reliqui do-I have given the dearly beloved of my soul into the hand of her enemies.

mum meam dimisi hereditatem meam dedi dilectam animam

12:8 Mine heritage is unto me as a lion in the forest; it meam in manu inimicorum eius 12:8 facta est mihi heredi-

crieth out against me: therefore have I hated it.

tas mea quasi leo in silua dedit contra me uocem ideo odiui

12:9 Mine heritage is unto me as a speckled bird, the birds eam 12:9 numquid auis discolor hereditas mea mihi numround about are against her; come ye, assemble all the beasts of the field, come to devour.

quid auis tincta per totum uenite congregamini omnes be-

12:10 Many pastors have destroyed my vineyard, they stiae terrae properate ad deuorandum 12:10 pastores multi

have trodden my portion under foot, they have made my pleasant portion a desolate wilderness.

demoliti sunt uineam meam conculcauerunt partem meam dederunt portionem meam desiderabilem in desertum soli-

12:11 They have made it desolate, and being desolate it tudinis 12:11 posuerunt eam in dissipationem luxitque su-

mourneth unto me; the whole land is made desolate, because no man layeth it to heart.

per me desolatione desolata est omnis terra quia nullus est

12:12 The spoilers are come upon all high places through qui recogitet corde 12:12 super omnes uias deserti uenerunt

the wilderness: for the sword of the LORD shall devour from the one end of the land even to the other end of the

land: no flesh shall have peace. uastatores quia gladius Domini deuorauit ab extremo terrae 12:13 They have sown wheat, but shall reap thorns: they usque ad extremum eius non est pax uniuersae carni 12:13

have put themselves to pain, but shall not profit: and they shall be ashamed of your revenues because of the fierce

anger of the LORD. seminauerunt triticum et spinas messuerunt hereditatem ac-

ceperunt et non eis prodeeit confundemini a fructibus uestris

12:14 Thus saith the LORD against all mine evil neigh- propter iram furoris Domini 12:14 haec dicit Dominus ad-

bours, that touch the inheritance which I have caused my people Israel to inherit; Behold, I will pluck them out of their land, and pluck out the house of Judah from among them.

uersum omnes uicinos meos pessimos qui tangunt hereditatem quam distribui populo meo Israhel ecce ego euellam eos de terra eorum et domum Iuda euellam de medio eorum

12:15 And it shall come to pass, after that I have plucked 12:15 et cum euellero eos conuertar et miserebor eorum et

them out I will return, and have compassion on them, and will bring them again, every man to his heritage, and every

man to his land. reducam eos uirum ad hereditatem suam et uirum in terram

Prophetia Ieremiae

suam 12:16 et erit si eruditi didicerint uias populi mei ut 12:16 And it shall come to pass, if they will diligently

learn the ways of my people, to swear by my name, The LORD liveth; as they taught my people to swear by Baal;

iurent in nomine meo uiuit Dominus sicut docuerunt popu- then shall they be built in the midst of my people. lum meum iurare in Baal aedificabuntur in medio populi mei 12:17 quod si non audierint euellam gentem illam euulsione 12:17 But if they will not obey, I will utterly pluck up and

destroy that nation, saith the LORD.

et perditione ait Dominus

13:1 haec dicit Dominus ad me uade et posside tibi lum- 13:1 Thus saith the LORD unto me, Go and get thee a

linen girdle, and put it upon thy loins, and put it not in water.

bare lineum et pones illud super lumbos tuos et in aquam non inferes illud 13:2 et possedi lumbare iuxta uerbum Do- 13:2 So I got a girdle according to the word of the LORD,

and put it on my loins.

mini et posui circa lumbos meos 13:3 et factus est sermo 13:3 And the word of the LORD came unto me the second

time, saying,

Domini ad me secundo dicens 13:4 tolle lumbare quod pos- 13:4 Take the girdle that thou hast got, which is upon thy

loins, and arise, go to Euphrates, and hide it there in a hole of the rock.

sedisti quod est circa lumbos tuos et surgens uade ad Eufraten et absconde illud ibi in foramine petrae 13:5 et abii et 13:5 So I went, and hid it by Euphrates, as the LORD

commanded me.

abscondi illud in Eufraten sicut praeceperat mihi Dominus 13:6 et factum est post dies plurimos dixit Dominus ad me 13:6 And it came to pass after many days, that the LORD

said unto me, Arise, go to Euphrates, and take the girdle from thence, which I commanded thee to hide there.

surge uade ad Eufraten et tolle inde lumbare quod praecepi tibi ut absconderes illud ibi 13:7 et abii ad Eufraten et fodi 13:7 Then I went to Euphrates, and digged, and took the

girdle from the place where I had hid it: and, behold, the girdle was marred, it was profitable for nothing.

et tuli lumbare de loco ubi absconderam illud et ecce conputruerat lumbare ita ut nullo usui aptum esset 13:8 et factum 13:8 Then the word of the LORD came unto me, saying, est uerbum Domini ad me dicens 13:9 haec dicit Dominus 13:9 Thus saith the LORD, After this manner will I mar

the pride of Judah, and the great pride of Jerusalem.

sic putrescere faciam superbiam Iuda et superbiam Hierusalem multam 13:10 populum istum pessimum qui nolunt 13:10 This evil people, which refuse to hear my words,

which walk in the imagination of their heart, and walk after other gods, to serve them, and to worship them, shall

audire uerba mea et ambulant in prauitate cordis sui abie- even be as this girdle, which is good for nothing. runtque post deos alienos ut seruirent eis et adorarent eos et erunt sicut lumbare istud quod nullo usui aptum est 13:11 13:11 For as the girdle cleaveth to the loins of a man, so

have I caused to cleave unto me the whole house of Israel and the whole house of Judah, saith the LORD; that they might be unto me for a people, and for a name, and for a praise, and for a glory: but they would not hear.

sicut enim adheret lumbare ad lumbos uiri sic adglutinaui mihi omnem domum Israhel et omnem domum Iuda dicit Dominus ut esset mihi in populum et in nomen et in laudem et in gloriam et non audierunt 13:12 dices ergo ad eos 13:12 Therefore thou shalt speak unto them this word;

Thus saith the LORD God of Israel, Every bottle shall be filled with wine: and they shall say unto thee, Do we not

sermonem istum haec dicit Dominus Deus Israhel omnis la- certainly know that every bottle shall be filled with wine?

guncula implebitur uino et dicent ad te numquid ignoramus quia omnis laguncula implebitur uino 13:13 et dices ad eos 13:13 Then shalt thou say unto them, Thus saith the

LORD, Behold, I will fill all the inhabitants of this land, even the kings that sit upon David’s throne, and the priests, and the prophets, and all the inhabitants of Jerusalem, with drunkenness.

haec dicit Dominus ecce ego implebo omnes habitatores terrae huius et reges qui sedent de stirpe Dauid super thronum

Biblia Sacra Vulgata

eius et sacerdotes et prophetas et omnes habitatores Hieru-

13:14 And I will dash them one against another, even the salem ebrietate 13:14 et dispergam eos uirum a fratre suo et

fathers and the sons together, saith the LORD: I will not pity, nor spare, nor have mercy, but destroy them.

patres et filios pariter ait Dominus non parcam et non con-

13:15 Hear ye, and give ear; be not proud: for the LORD cedam neque miserebor ut non disperdam eos 13:15 audite

hath spoken.

et auribus percipite nolite eleuari quia Dominus locutus est

13:16 Give glory to the LORD your God, before he cause 13:16 date Domino Deo uestro gloriam antequam contene-

darkness, and before your feet stumble upon the dark mountains, and, while ye look for light, he turn it into the

shadow of death, and make it gross darkness. brescat et antequam offendant pedes uestri ad montes cali-

gosos expectabitis lucem et ponet eam in umbram mortis et

13:17 But if ye will not hear it, my soul shall weep in se- in caliginem 13:17 quod si hoc non audieritis in abscondito

cret places for your pride; and mine eye shall weep sore, and run down with tears, because the LORD’s flock is car-

ried away captive. plorabit anima mea a facie superbiae plorans plorabit et de-

ducet oculus meus lacrimam quia captus est grex Domini

13:18 Say unto the king and to the queen, Humble your- 13:18 dic regi et dominatrici humiliamini sedete quoniam de-

selves, sit down: for your principalities shall come down, even the crown of your glory.

scendit de capite uestro corona gloriae uestrae 13:19 ciui-

13:19 The cities of the south shall be shut up, and none shall open them: Judah shall be carried away captive all of it, it shall be wholly carried away captive.

tates austri clausae sunt et non est qui aperiat translata est

13:20 Lift up your eyes, and behold them that come from omnis Iudaea transmigratione perfecta 13:20 leuate oculos

the north: where is the flock that was given thee, thy beautiful flock?

uestros et uidete qui uenitis ab aquilone ubi est grex qui da-

13:21 What wilt thou say when he shall punish thee? for tus est tibi pecus inclitum tuum 13:21 quid dices cum uisi-

thou hast taught them to be captains, and as chief over thee: shall not sorrows take thee, as a woman in travail?

tauerit te tu enim docuisti eos aduersum te et erudisti in caput tuum numquid non dolores adprehendent te quasi mulierem

13:22 And if thou say in thine heart, Wherefore come parturientem 13:22 quod si dixeris in corde tuo quare ue-

these things upon me? For the greatness of thine iniquity are thy skirts discovered, and thy heels made bare.

nerunt mihi haec propter multitudinem iniquitatis tuae reue-

13:23 Can the Ethiopian change his skin, or the leopard lata sunt uerecundiora tua pollutae sunt plantae tuae 13:23

his spots? then may ye also do good, that are accustomed to do evil.

si mutare potest Aethiops pellem suam aut pardus uarietates suas et uos poteritis bene facere cum didiceritis malum

13:24 Therefore will I scatter them as the stubble that pas- 13:24 et disseminabo eos quasi stipulam quae uento raptatur

seth away by the wind of the wilderness.

13:25 This is thy lot, the portion of thy measures from in deserto 13:25 haec sors tua parsque mensurae tuae a me

me, saith the LORD; because thou hast forgotten me, and trusted in falsehood.

dicit Dominus quia oblita es mei et confisa es in mendacio

13:26 Therefore will I discover thy skirts upon thy face, 13:26 unde et ego nudaui femora tua contra faciem tuam et that thy shame may appear.

13:27 I have seen thine adulteries, and thy neighings, the apparuit ignominia tua 13:27 adulteria tua et hinnitus tuus

lewdness of thy whoredom, and thine abominations on the hills in the fields. Woe unto thee, O Jerusalem! wilt thou

not be made clean? when shall it once be? scelus fornicationis tuae super colles in agro uidi abominationes tuas uae tibi Hierusalem non mundaberis post me usquequo adhuc

14:1 The word of the LORD that came to Jeremiah con- 14:1 quod factum est uerbum Domini ad Hieremiam de cerning the dearth.

Prophetia Ieremiae

sermonibus siccitatis 14:2 luxit Iudaea et portae eius cor- 14:2 Judah mourneth, and the gates thereof languish; they

are black unto the ground; and the cry of Jerusalem is gone up.

ruerunt et obscuratae sunt in terra et clamor Hierusalem as-

cendit 14:3 maiores miserunt minores suos ad aquam ue- 14:3 And their nobles have sent their little ones to the waters: they came to the pits, and found no water; they returned with their vessels empty; they were ashamed and

nerunt ad hauriendum non inuenerunt aquam reportauerunt confounded, and covered their heads. uasa sua uacua confusi sunt et adflicti et operuerunt capita sua 14:4 propter terrae uastitatem quia non uenit pluuia in 14:4 Because the ground is chapt, for there was no rain in

the earth, the plowmen were ashamed, they covered their heads.

terra confusi sunt agricolae operuerunt capita sua 14:5 nam

14:5 Yea, the hind also calved in the field, and forsook it, because there was no grass.

et cerua in agro peperit et reliquit quia non erat herba 14:6 14:6 And the wild asses did stand in the high places, they

snuffed up the wind like dragons; their eyes did fail, because there was no grass.

et onagri steterunt in rupibus traxerunt uentum quasi dra-

cones defecerunt oculi eorum quia non erat herba 14:7 si 14:7 O LORD, though our iniquities testify against us, do thou it for thy name’s sake: for our backslidings are many; we have sinned against thee.

iniquitates nostrae responderunt nobis Domine fac propter nomen tuum quoniam multae sunt auersiones nostrae tibi peccauimus 14:8 expectatio Israhel saluator eius in tem- 14:8 O the hope of Israel, the saviour thereof in time of

trouble, why shouldest thou be as a stranger in the land, and as a wayfaring man that turneth aside to tarry for a

pore tribulationis quare quasi colonus futurus es in terra et night? quasi uiator declinans ad manendum 14:9 quare futurus es 14:9 Why shouldest thou be as a man astonied, as a mighty

man that cannot save? yet thou, O LORD, art in the midst of us, and we are called by thy name; leave us not.

uelut uir uagus ut fortis qui non potest saluare tu autem in nobis es Domine et nomen tuum super nos inuocatum est ne derelinquas nos 14:10 haec dicit Dominus populo huic qui 14:10 Thus saith the LORD unto this people, Thus have

they loved to wander, they have not refrained their feet, therefore the LORD doth not accept them; he will now

dilexit mouere pedes suos et non quieuit et Domino non pla- remember their iniquity, and visit their sins. cuit nunc recordabitur iniquitatum eorum et uisitabit peccata eorum 14:11 et dixit Dominus ad me noli orare pro populo 14:11 Then said the LORD unto me, Pray not for this peo-

ple for their good.

isto in bonum 14:12 cum ieiunauerint non exaudiam pre- 14:12 When they fast, I will not hear their cry; and when

they offer burnt offering and an oblation, I will not accept them: but I will consume them by the sword, and by the

ces eorum et si obtulerint holocaustomata et uictimas non famine, and by the pestilence. suscipiam ea quoniam gladio et fame et peste ego consumam eos 14:13 et dixi a a a Domine Deus prophetae dicunt 14:13 Then said I, Ah, Lord GOD! behold, the prophets

say unto them, Ye shall not see the sword, neither shall ye have famine; but I will give you assured peace in this

eis non uidebitis gladium et famis non erit in uobis sed pa- place. cem ueram dabit uobis in loco isto 14:14 et dixit Dominus 14:14 Then the LORD said unto me, The prophets proph-

esy lies in my name: I sent them not, neither have I commanded them, neither spake unto them: they prophesy unto you a false vision and divination, and a thing of nought, and the deceit of their heart.

ad me falso prophetae uaticinantur in nomine meo non misi eos et non praecepi eis neque locutus sum ad eos uisionem mendacem et diuinationem et fraudulentiam et seductionem cordis sui prophetant uobis 14:15 ideo haec dicit Dominus 14:15 Therefore thus saith the LORD concerning the

prophets that prophesy in my name, and I sent them not, yet they say, Sword and famine shall not be in this land;

de prophetis qui prophetant in nomine meo quos ego non By sword and famine shall those prophets be consumed. misi dicentes gladius et famis non erit in terra hac in gla-

Biblia Sacra Vulgata

14:16 And the people to whom they prophesy shall be cast dio et fame consumentur prophetae illi 14:16 et populi qui-

out in the streets of Jerusalem because of the famine and the sword; and they shall have none to bury them, them, their wives, nor their sons, nor their daughters: for I will pour their wickedness upon them.

bus prophetant erunt proiecti in uiis Hierusalem prae fame et gladio et non erit qui sepeliat eos ipsi et uxores eorum filii

14:17 Therefore thou shalt say this word unto them; Let et filiae eorum et effundam super eos malum suum 14:17 et mine eyes run down with tears night and day, and let them not cease: for the virgin daughter of my people is broken

with a great breach, with a very grievous blow. dices ad eos uerbum istud deducant oculi mei lacrimam per noctem et diem et non taceant quoniam contritione magna contrita est uirgo filia populi mei plaga pessima uehemen-

14:18 If I go forth into the field, then behold the slain with ter 14:18 si egressus fuero ad agros ecce occisi gladio et si

the sword! and if I enter into the city, then behold them that are sick with famine! yea, both the prophet and the

priest go about into a land that they know not. introiero in ciuitatem ecce adtenuati fame propheta quoque 14:19 Hast thou utterly rejected Judah? hath thy soul et sacerdos abierunt in terram quam ignorabant 14:19 num-

lothed Zion? why hast thou smitten us, and there is no healing for us? we looked for peace, and there is no good;

and for the time of healing, and behold trouble! quid proiciens abiecisti Iudam aut Sion abominata est anima tua quare ergo percussisti nos ita ut nulla sit sanitas expectauimus pacem et non est bonum et tempus curationis et ecce

14:20 We acknowledge, O LORD, our wickedness, and turbatio 14:20 cognouimus Domine impietates nostras in-

the iniquity of our fathers: for we have sinned against thee.

14:21 Do not abhor us, for thy name’s sake, do not dis- iquitatem patrum nostrorum quia peccauimus tibi 14:21 ne

grace the throne of thy glory: remember, break not thy covenant with us.

nos des in obprobrium propter nomen tuum neque facias nobis contumeliam solii gloriae tuae recordare ne irritum facias

14:22 Are there any among the vanities of the Gentiles that foedus tuum nobiscum 14:22 numquid sunt in sculptilibus can cause rain? or can the heavens give showers? art not thou he, O LORD our God? therefore we will wait upon

thee: for thou hast made all these things. gentium qui pluant aut caeli possunt dare imbres nonne tu es Domine Deus noster quem expectauimus tu enim fecisti omnia haec

15:1 Then said the LORD unto me, Though Moses and 15:1 et dixit Dominus ad me si steterit Moses et Samuhel

Samuel stood before me, yet my mind could not be toward this people: cast them out of my sight, and let them go

forth. coram me non est anima mea ad populum istum eice illos 15:2 And it shall come to pass, if they say unto thee, a facie mea et egrediantur 15:2 quod si dixerint ad te quo

Whither shall we go forth? then thou shalt tell them, Thus saith the LORD; Such as are for death, to death; and such as are for the sword, to the sword; and such as are for the famine, to the famine; and such as are for the captivity, to the captivity.

egrediemur dices ad eos haec dicit Dominus qui ad mortem ad mortem et qui ad gladium ad gladium et qui ad famem

15:3 And I will appoint over them four kinds, saith the ad famem et qui ad captiuitatem ad captiuitatem 15:3 et ui-

LORD: the sword to slay, and the dogs to tear, and the fowls of the heaven, and the beasts of the earth, to devour

and destroy. sitabo super eos quattuor species dicit Dominus gladium ad

occisionem et canes ad lacerandum et uolatilia caeli et bes-

15:4 And I will cause them to be removed into all tias terrae ad deuorandum et dissipandum 15:4 et dabo eos

kingdoms of the earth, because of Manasseh the son of Hezekiah king of Judah, for that which he did in

Jerusalem. in feruorem uniuersis regnis terrae propter Manassem filium

Ezechiae regis Iuda super omnibus quae fecit in Hierusalem

15:5 For who shall have pity upon thee, O Jerusalem? or 15:5 quis enim miserebitur tui Hierusalem aut quis contrista-

who shall bemoan thee? or who shall go aside to ask how thou doest?

Prophetia Ieremiae

bitur pro te aut quis ibit ad rogandum pro pace tua 15:6 tu 15:6 Thou hast forsaken me, saith the LORD, thou art

gone backward: therefore will I stretch out my hand against thee, and destroy thee; I am weary with repenting.

reliquisti me dicit Dominus retrorsum abisti et extendam ma-

num meam super te et interficiam te laboraui rogans 15:7 15:7 And I will fan them with a fan in the gates of the land; I will bereave them of children, I will destroy my people since they return not from their ways.

et dispergam eos uentilabro in portis terrae interfeci et perdidi populum meum et tamen a uiis suis non sunt reuersi 15:8 multiplicatae sunt mihi uiduae eius super harenam ma- 15:8 Their widows are increased to me above the sand of

the seas: I have brought upon them against the mother of the young men a spoiler at noonday: I have caused him to

ris induxi eis super matrem adulescentis uastatorem meri- fall upon it suddenly, and terrors upon the city. die misi super ciuitates repente terrorem 15:9 infirmata est 15:9 She that hath borne seven languisheth: she hath given

up the ghost; her sun is gone down while it was yet day: she hath been ashamed and confounded: and the residue of them will I deliver to the sword before their enemies, saith the LORD.

quae peperit septem defecit anima eius occidit ei sol cum adhuc esset dies confusa est et erubuit et residuos eius in gladium dabo in conspectu inimicorum eorum ait Dominus 15:10 uae mihi mater mea quare genuisti me uirum rixae ui- 15:10 Woe is me, my mother, that thou hast borne me a

man of strife and a man of contention to the whole earth! I have neither lent on usury, nor men have lent to me on

rum discordiae in uniuersa terra non feneraui nec fenerauit usury; yet every one of them doth curse me.

mihi quisquam omnes maledicunt mihi 15:11 dicit Domi- 15:11 The LORD said, Verily it shall be well with thy remnant; verily I will cause the enemy to entreat thee well in the time of evil and in the time of affliction.

nus si non reliquiae tuae in bonum si non occurri tibi in tempore adflictionis et in tempore tribulationis aduersum inimicum 15:12 numquid foederabitur ferrum ferro ab aquilone 15:12 Shall iron break the northern iron and the steel? et aes 15:13 diuitias tuas et thesauros tuos in direptionem 15:13 Thy substance and thy treasures will I give to the

spoil without price, and that for all thy sins, even in all thy borders.

dabo gratis in omnibus peccatis tuis et in omnibus terminis tuis 15:14 et adducam inimicos tuos de terra qua nes- 15:14 And I will make thee to pass with thine enemies into

a land which thou knowest not: for a fire is kindled in mine anger, which shall burn upon you.

cis quia ignis succensus est in furore meo super uos ardebit 15:15 tu scis Domine recordare mei et uisita me et tuere me 15:15 O LORD, thou knowest: remember me, and visit

me, and revenge me of my persecutors; take me not away in thy longsuffering: know that for thy sake I have suffered

ab his qui persequuntur me noli in patientia tua suscipere rebuke. me scito quoniam sustinui pro te obprobrium 15:16 inuenti 15:16 Thy words were found, and I did eat them; and thy

word was unto me the joy and rejoicing of mine heart: for I am called by thy name, O LORD God of hosts.

sunt sermones tui et comedi eos et factum est mihi uerbum tuum in gaudium et in laetitiam cordis mei quoniam inuocatum est nomen tuum super me Domine Deus exercituum 15:17 non sedi in concilio ludentium et gloriatus sum a facie 15:17 I sat not in the assembly of the mockers, nor re-

joiced; I sat alone because of thy hand: for thou hast filled me with indignation.

manus tuae solus sedebam quoniam comminatione replesti me 15:18 quare factus est dolor meus perpetuus et plaga 15:18 Why is my pain perpetual, and my wound incurable,

which refuseth to be healed? wilt thou be altogether unto me as a liar, and as waters that fail?

mea desperabilis rennuit curari facta est mihi quasi menda-

cium aquarum infidelium 15:19 propter hoc haec dicit Do- 15:19 Therefore thus saith the LORD, If thou return, then will I bring thee again, and thou shalt stand before me: and if thou take forth the precious from the vile, thou shalt be as my mouth: let them return unto thee; but return not thou unto them.

minus si conuerteris conuertam te et ante faciem meam sta-

Biblia Sacra Vulgata

bis et si separaueris pretiosum a uili quasi os meum eris con-

15:20 And I will make thee unto this people a fenced uertentur ipsi ad te et tu non conuerteris ad eos 15:20 et

brasen wall: and they shall fight against thee, but they shall not prevail against thee: for I am with thee to save

thee and to deliver thee, saith the LORD. dabo te populo huic in murum aereum fortem et bellabunt

aduersum te et non praeualebunt quia ego tecum sum ut sal-

15:21 And I will deliver thee out of the hand of the wicked, uem te et eruam dicit Dominus 15:21 et liberabo te de manu and I will redeem thee out of the hand of the terrible.

pessimorum et rediiam te de manu fortium

16:1 The word of the LORD came also unto me, saying, 16:1 et factum est uerbum Domini ad me dicens 16:2 non

16:2 Thou shalt not take thee a wife, neither shalt thou have sons or daughters in this place.

accipies uxorem et non erunt tibi filii et filiae in loco isto

16:3 For thus saith the LORD concerning the sons and 16:3 quia haec dicit Dominus super filios et filias qui gene-

concerning the daughters that are born in this place, and concerning their mothers that bare them, and concerning

their fathers that begat them in this land; rantur in loco isto et super matres eorum quae genuerunt eos

et super patres eorum de quorum stirpe sunt nati in terra

16:4 They shall die of grievous deaths; they shall not be hac 16:4 mortibus aegrotationum morientur non plangen-

lamented; neither shall they be buried; but they shall be as dung upon the face of the earth: and they shall be consumed by the sword, and by famine; and their carcases shall be meat for the fowls of heaven, and for the beasts of the earth.

tur et non sepelientur in sterquilinium super faciem terrae erunt et gladio et fame consumentur et erit cadauer eorum in

16:5 For thus saith the LORD, Enter not into the house escam uolatilibus caeli et bestiis terrae 16:5 haec enim di-

of mourning, neither go to lament nor bemoan them: for I have taken away my peace from this people, saith the

LORD, even lovingkindness and mercies. cit Dominus ne ingrediaris domum conuiuii neque uadas ad plangendum neque consoleris eos quia abstuli pacem meam a populo isto dicit Dominus misericordiam et miserationes

16:6 Both the great and the small shall die in this land: 16:6 et morientur grandes et parui in terra ista non sepelientur

they shall not be buried, neither shall men lament for them, nor cut themselves, nor make themselves bald for them:

neque plangentur et non se incident neque caluitium fiet pro

16:7 Neither shall men tear themselves for them in mourn- eis 16:7 et non frangent inter eos lugenti panem ad consoing, to comfort them for the dead; neither shall men give them the cup of consolation to drink for their father or for

their mother. landum super mortuo et non dabunt eis potum calicis ad con-

16:8 Thou shalt not also go into the house of feasting, to solandum super patre suo et matre 16:8 et domum conuiuii sit with them to eat and to drink.

16:9 For thus saith the LORD of hosts, the God of Israel; non ingredieris ut sedeas cum eis et comedas et bibas 16:9

Behold, I will cause to cease out of this place in your eyes, and in your days, the voice of mirth, and the voice of gladness, the voice of the bridegroom, and the voice of the bride.

quia haec dicit Dominus exercituum Deus Israhel ecce ego auferam de loco isto in oculis uestris et in diebus uestris uocem gaudii et uocem laetitiae uocem sponsi et uocem spon-

16:10 And it shall come to pass, when thou shalt shew sae 16:10 et cum adnuntiaueris populo huic omnia uerba

this people all these words, and they shall say unto thee, Wherefore hath the LORD pronounced all this great evil against us? or what is our iniquity? or what is our sin that we have committed against the LORD our God?

haec et dixerint tibi quare locutus est Dominus super nos omne malum grande istud quae iniquitas nostra et quod pec-

16:11 Then shalt thou say unto them, Because your fathers catum nostrum quod peccauimus Domino Deo nostro 16:11 have forsaken me, saith the LORD, and have walked after other gods, and have served them, and have worshipped

them, and have forsaken me, and have not kept my law; dices ad eos quia dereliquerunt patres uestri me ait Dominus et abierunt post deos alienos et seruierunt eis et adorauerunt

Prophetia Ieremiae

eos et me dereliquerunt et legem meam non custodierunt 16:12 sed et uos peius operati estis quam patres uestri ecce 16:12 And ye have done worse than your fathers; for, be-

hold, ye walk every one after the imagination of his evil heart, that they may not hearken unto me:

enim ambulat unusquisque post prauitatem cordis sui mali

ut me non audiat 16:13 et eiciam uos de terra hac in terram 16:13 Therefore will I cast you out of this land into a land that ye know not, neither ye nor your fathers; and there shall ye serve other gods day and night; where I will not

quam ignoratis uos et patres uestri et seruietis ibi diis alienis shew you favour.

die ac nocte qui non dabunt uobis requiem 16:14 propte- 16:14 Therefore, behold, the days come, saith the LORD, that it shall no more be said, The LORD liveth, that brought up the children of Israel out of the land of Egypt;

rea ecce dies ueniunt dicit Dominus et non dicetur ultra uiuit Dominus qui eduxit filios Israhel de terra Aegypti 16:15 16:15 But, The LORD liveth, that brought up the children

of Israel from the land of the north, and from all the lands whither he had driven them: and I will bring them again

sed uiuit Dominus qui eduxit filios Israhel de terra aquilo- into their land that I gave unto their fathers. nis et de uniuersis terris ad quas eieci eos et reducam eos in terram suam quam dedi patribus eorum 16:16 ecce ego mit- 16:16 Behold, I will send for many fishers, saith the

LORD, and they shall fish them; and after will I send for many hunters, and they shall hunt them from every moun-

tam piscatores multos dicit Dominus et piscabuntur eos et tain, and from every hill, and out of the holes of the rocks. post haec mittam eis multos uenatores et uenabuntur eos de omni monte et de omni colle et de cauernis petrarum 16:17 16:17 For mine eyes are upon all their ways: they are not

hid from my face, neither is their iniquity hid from mine eyes.

quia oculi mei super omnes uias eorum non sunt absconditae a facie mea et non fuit occulta iniquitas eorum ab oculis meis 16:18 et reddam primum duplices iniquitates et pec- 16:18 And first I will recompense their iniquity and their

sin double; because they have defiled my land, they have filled mine inheritance with the carcases of their detestable

cata eorum quia contaminauerunt terram meam in morticinis and abominable things. idolorum suorum et abominationibus suis impleuerunt hereditatem meam 16:19 Domine fortitudo mea et robur meum 16:19 O LORD, my strength, and my fortress, and my

refuge in the day of affliction, the Gentiles shall come unto thee from the ends of the earth, and shall say, Surely our fathers have inherited lies, vanity, and things wherein there is no profit.

et refugium meum in die tribulationis ad te gentes uenient ab extremis terrae et dicent uere mendacium possederunt patres nostri uanitatem quae eis non profuit 16:20 numquid faciet 16:20 Shall a man make gods unto himself, and they are

no gods?

sibi homo deos et ipsi non sunt dii 16:21 idcirco ecce ego 16:21 Therefore, behold, I will this once cause them to

know, I will cause them to know mine hand and my might; and they shall know that my name is The LORD.

ostendam eis per uicem hanc ostendam eis manum meam et uirtutem meam et scient quia nomen mihi Dominus

17:1 peccatum Iuda scriptum est stilo ferreo in ungue ad- 17:1 The sin of Judah is written with a pen of iron, and

with the point of a diamond: it is graven upon the table of their heart, and upon the horns of your altars;

amantino exaratum super latitudinem cordis eorum et in

cornibus ararum eorum 17:2 cum recordati fuerint filii eo- 17:2 Whilst their children remember their altars and their

groves by the green trees upon the high hills.

rum ararum suarum et lucorum lignorumque frondentium in montibus excelsis 17:3 sacrificantes in agro fortitudinem 17:3 O my mountain in the field, I will give thy substance

and all thy treasures to the spoil, and thy high places for sin, throughout all thy borders.

tuam et omnes thesauros tuos in direptionem dabo excelsa tua propter peccata in uniuersis finibus tuis 17:4 et relin- 17:4 And thou, even thyself, shalt discontinue from thine

heritage that I gave thee; and I will cause thee to serve thine enemies in the land which thou knowest not: for ye have kindled a fire in mine anger, which shall burn for ever.

Biblia Sacra Vulgata

queris sola ab hereditate tua quam dedi tibi et seruire te faciam inimicis tuis in terra quam ignoras quoniam ignem suc-

17:5 Thus saith the LORD; Cursed be the man that trusteth cendisti in furore meo usque in aeternum ardebit 17:5 haec

in man, and maketh flesh his arm, and whose heart departeth from the LORD.

dicit Dominus maledictus homo qui confidit in homine et ponit carnem brachium suum et a Domino recedit cor eius

17:6 For he shall be like the heath in the desert, and shall 17:6 erit enim quasi myrice in deserto et non uidebit cum not see when good cometh; but shall inhabit the parched places in the wilderness, in a salt land and not inhabited.

uenerit bonum sed habitabit in siccitate in deserto in terra

17:7 Blessed is the man that trusteth in the LORD, and salsuginis et inhabitabili 17:7 benedictus uir qui confidit in

whose hope the LORD is.

17:8 For he shall be as a tree planted by the waters, and Domino et erit Dominus fiducia eius 17:8 et erit quasi lig-

that spreadeth out her roots by the river, and shall not see when heat cometh, but her leaf shall be green; and shall not be careful in the year of drought, neither shall cease from yielding fruit.

num quod transplantatur super aquas quod ad humorem mittit radices suas et non timebit cum uenerit aestus et erit folium eius uiride et in tempore siccitatis non erit sollicitum

17:9 The heart is deceitful above all things, and desper- nec aliquando desinet facere fructum 17:9 prauum est cor ately wicked: who can know it?

17:10 I the LORD search the heart, I try the reins, even omnium et inscrutabile quis cognoscet illud 17:10 ego Do-

to give every man according to his ways, and according to the fruit of his doings.

minus scrutans cor et probans renes qui do unicuique iuxta

17:11 As the partridge sitteth on eggs, and hatcheth them uiam et iuxta fructum adinuentionum suarum 17:11 perdix not; so he that getteth riches, and not by right, shall leave them in the midst of his days, and at his end shall be a fool.

fouit quae non peperit fecit diuitias et non in iudicio in dimidio dierum suorum derelinquet eas et in nouissimo suo

17:12 A glorious high throne from the beginning is the erit insipiens 17:12 solium gloriae altitudinis a principio

place of our sanctuary.

17:13 O LORD, the hope of Israel, all that forsake thee locus sanctificationis nostrae 17:13 expectatio Israhel Do-

shall be ashamed, and they that depart from me shall be written in the earth, because they have forsaken the

LORD, the fountain of living waters. mine omnes qui te derelinquunt confundentur recedentes in

terra scribentur quoniam dereliquerunt uenam aquarum ui-

17:14 Heal me, O LORD, and I shall be healed; save me, uentium Dominum 17:14 sana me Domine et sanabor sal-

and I shall be saved: for thou art my praise. 17:15 Behold, they say unto me, Where is the word of the uum me fac et saluus ero quoniam laus mea tu es 17:15 LORD? let it come now.

17:16 As for me, I have not hastened from being a pastor ecce ipsi dicunt ad me ubi est uerbum Domini ueniat 17:16

to follow thee: neither have I desired the woeful day; thou knowest: that which came out of my lips was right before

thee. et ego non sum turbatus te pastorem sequens et diem homi-

nis non desideraui tu scis quod egressum est de labiis meis

17:17 Be not a terror unto me: thou art my hope in the day rectum in conspectu tuo fuit 17:17 non sis mihi tu formi-

of evil.

17:18 Let them be confounded that persecute me, but let dini spes mea tu in die adflictionis 17:18 confundantur qui

not me be confounded: let them be dismayed, but let not me be dismayed: bring upon them the day of evil, and

destroy them with double destruction. persequuntur me et non confundar ego paueant illi et non pa-

ueam ego induc super eos diem adflictionis et duplici con-

17:19 Thus said the LORD unto me; Go and stand in the tritione contere eos 17:19 haec dicit Dominus ad me uade

gate of the children of the people, whereby the kings of Judah come in, and by the which they go out, and in all

the gates of Jerusalem; et sta in porta filiorum populi per quam ingrediuntur reges

Prophetia Ieremiae

Iuda et egrediuntur et in cunctis portis Hierusalem 17:20 et 17:20 And say unto them, Hear ye the word of the LORD,

ye kings of Judah, and all Judah, and all the inhabitants of Jerusalem, that enter in by these gates:

dices ad eos audite uerbum Domini reges Iuda et omnis Iudaea cunctique habitatores Hierusalem qui ingredimini per portas istas 17:21 haec dicit Dominus custodite animas ue- 17:21 Thus saith the LORD; Take heed to yourselves, and

bear no burden on the sabbath day, nor bring it in by the gates of Jerusalem;

stras et nolite portare pondera in die sabbati nec inferatis per portas Hierusalem 17:22 et nolite eicere onera de domi- 17:22 Neither carry forth a burden out of your houses on

the sabbath day, neither do ye any work, but hallow ye the sabbath day, as I commanded your fathers.

bus uestris in die sabbati et omne opus non facietis sanctificate diem sabbati sicut praecepi patribus uestris 17:23 et 17:23 But they obeyed not, neither inclined their ear, but

made their neck stiff, that they might not hear, nor receive instruction.

non audierunt nec inclinauerunt aurem suam sed indurauerunt ceruicem suam ne audirent me et ne acciperent disciplinam 17:24 et erit si audieritis me dicit Dominus ut non 17:24 And it shall come to pass, if ye diligently hearken

unto me, saith the LORD, to bring in no burden through the gates of this city on the sabbath day, but hallow the

inferatis onera per portas ciuitatis huius in die sabbati et si sabbath day, to do no work therein; sanctificaueritis diem sabbati ne faciatis in ea omne opus 17:25 ingredientur per portas ciuitatis huius reges et princi- 17:25 Then shall there enter into the gates of this city kings

and princes sitting upon the throne of David, riding in chariots and on horses, they, and their princes, the men of Judah, and the inhabitants of Jerusalem: and this city shall remain for ever.

pes sedentes super solium Dauid et ascendentes in curribus et equis ipsi et principes eorum uir Iuda et habitatores Hierusalem et habitabitur ciuitas haec in sempiternum 17:26 et 17:26 And they shall come from the cities of Judah, and

from the places about Jerusalem, and from the land of Benjamin, and from the plain, and from the mountains, and from the south, bringing burnt offerings, and sacrifices, and meat offerings, and incense, and bringing sacrifices of praise, unto the house of the LORD.

uenient de ciuitate Iuda et de circuitu Hierusalem et de terra Beniamin et de campestribus et de montuosis et ab austro portantes holocaustum et uictimam et sacrificium et tus et inferent oblationem in domum Domini 17:27 si autem non 17:27 But if ye will not hearken unto me to hallow the

sabbath day, and not to bear a burden, even entering in at the gates of Jerusalem on the sabbath day; then will I kindle a fire in the gates thereof, and it shall devour the palaces of Jerusalem, and it shall not be quenched.

audieritis me ut sanctificetis diem sabbati et ne portetis onus et ne inferatis per portas Hierusalem in die sabbati succendam ignem in portis eius et deuorabit domos Hierusalem et non extinguetur

18:1 uerbum quod factum est ad Hieremiam a Domino 18:1 The word which came to Jeremiah from the LORD,

saying,

dicens 18:2 surge et descende in domum figuli et ibi au- 18:2 Arise, and go down to the potter’s house, and there I

will cause thee to hear my words.

dies uerba mea 18:3 et descendi in domum figuli et ecce 18:3 Then I went down to the potter’s house, and, behold,

he wrought a work on the wheels.

ipse faciebat opus super rotam 18:4 et dissipatum est uas 18:4 And the vessel that he made of clay was marred in

the hand of the potter: so he made it again another vessel, as seemed good to the potter to make it.

quod ipse faciebat e luto manibus suis conuersusque fecit illud uas alterum sicut placuerat in oculis eius ut faceret 18:5 18:5 Then the word of the LORD came to me, saying,

et factum est uerbum Domini ad me dicens 18:6 numquid 18:6 O house of Israel, cannot I do with you as this potter? saith the LORD. Behold, as the clay is in the potter’s hand, so are ye in mine hand, O house of Israel.

sicut figulus iste non potero facere uobis domus Israhel ait

Biblia Sacra Vulgata

Dominus ecce sicut lutum in manu figuli sic uos in manu

18:7 At what instant I shall speak concerning a nation, and mea domus Israhel 18:7 repente loquar aduersum gentem concerning a kingdom, to pluck up, and to pull down, and to destroy it;

et aduersum regnum ut eradicem et destruam et disperdam

18:8 If that nation, against whom I have pronounced, turn illud 18:8 si paenitentiam egerit gens illa a malo suo quod

from their evil, I will repent of the evil that I thought to do unto them.

locutus sum aduersum eam agam et ego paenitentiam super

18:9 And at what instant I shall speak concerning a nation, malo quod cogitaui ut facerem ei 18:9 et subito loquar de and concerning a kingdom, to build and to plant it;

18:10 If it do evil in my sight, that it obey not my voice, gente et regno ut aedificem et ut plantem illud 18:10 si fe-

then I will repent of the good, wherewith I said I would benefit them.

cerit malum in oculis meis ut non audiat uocem meam paenitentiam agam super bono quod locutus sum ut facerem ei

18:11 Now therefore go to, speak to the men of Judah, 18:11 nunc ergo dic uiro Iudae et habitatoribus Hierusalem

and to the inhabitants of Jerusalem, saying, Thus saith the LORD; Behold, I frame evil against you, and devise a device against you: return ye now every one from his evil way, and make your ways and your doings good.

dicens haec dicit Dominus ecce ego fingo contra uos malum et cogito contra uos cogitationem reuertatur unusquisque a uia sua mala et dirigite uias uestras et studia uestra

18:12 And they said, There is no hope: but we will walk 18:12 qui dixerunt desperauimus post cogitationes enim no-

after our own devices, and we will every one do the imagination of his evil heart.

stras ibimus et unusquisque prauitatem cordis sui mali facie-

18:13 Therefore thus saith the LORD; Ask ye now among mus 18:13 ideo haec dicit Dominus interrogate gentes quis

the heathen, who hath heard such things: the virgin of Israel hath done a very horrible thing.

audiuit talia horribilia quae fecit nimis uirgo Israhel 18:14

18:14 Will a man leave the snow of Lebanon which cometh from the rock of the field? or shall the cold flowing waters that come from another place be forsaken?

numquid deficiet de petra agri nix Libani aut euelli possunt

18:15 Because my people hath forgotten me, they have aquae erumpentes frigidae et defluentes 18:15 quia oblitus

burned incense to vanity, and they have caused them to stumble in their ways from the ancient paths, to walk in

paths, in a way not cast up; est mei populus meus frustra libantes et inpingentes in uiis

suis in semitis saeculi ut ambularent per eas in itinere non

18:16 To make their land desolate, and a perpetual hiss- trito 18:16 ut fieret terra eorum in desolationem et in sibi-

ing; every one that passeth thereby shall be astonished, and wag his head.

lum sempiternum omnis qui praeterit per eam obstupescet

18:17 I will scatter them as with an east wind before the et mouebit caput suum 18:17 sicut uentus urens dispergam

enemy; I will shew them the back, and not the face, in the day of their calamity.

eos coram inimico dorsum et non faciem ostendam eis in

18:18 Then said they, Come and let us devise devices die perditionis eorum 18:18 et dixerunt uenite et cogitemus

against Jeremiah; for the law shall not perish from the priest, nor counsel from the wise, nor the word from the prophet. Come, and let us smite him with the tongue, and let us not give heed to any of his words.

contra Hieremiam cogitationes non enim peribit lex a sacerdote neque consilium a sapiente nec sermo a propheta uenite et percutiamus eum lingua et non adtendamus ad uniuersos

18:19 Give heed to me, O LORD, and hearken to the voice sermones eius 18:19 adtende Domine ad me et audi uocem

of them that contend with me. 18:20 Shall evil be recompensed for good? for they have aduersariorum meorum 18:20 numquid redditur pro bono digged a pit for my soul. Remember that I stood before thee to speak good for them, and to turn away thy wrath

from them. malum quia foderunt foueam animae meae recordare quod steterim in conspectu tuo ut loquerer pro eis bonum et auer-

Prophetia Ieremiae

terem indignationem tuam ab eis 18:21 propterea da filios 18:21 Therefore deliver up their children to the famine,

and pour out their blood by the force of the sword; and let their wives be bereaved of their children, and be widows; and let their men be put to death; let their young men be slain by the sword in battle.

eorum in famem et deduc eos in manus gladii fiant uxores eorum absque liberis et uiduae et uiri earum interficiantur morte iuuenes eorum confodiantur gladio in proelio 18:22 18:22 Let a cry be heard from their houses, when thou

shalt bring a troop suddenly upon them: for they have digged a pit to take me, and hid snares for my feet.

audiatur clamor de domibus eorum adduces enim super eos latronem repente quia foderunt foueam ut caperent me et la-

queos absconderunt pedibus meis 18:23 tu autem Domine 18:23 Yet, LORD, thou knowest all their counsel against me to slay me: forgive not their iniquity, neither blot out their sin from thy sight, but let them be overthrown before

scis omne consilium eorum aduersum me in mortem ne pro- thee; deal thus with them in the time of thine anger. pitieris iniquitati eorum et peccatum eorum a facie tua non deleatur fiant corruentes in conspectu tuo in tempore furoris tui abutere eis

19:1 haec dicit Dominus uade et accipe lagunculam figuli 19:1 Thus saith the LORD, Go and get a potter’s earthen

bottle, and take of the ancients of the people, and of the ancients of the priests;

testeam a senioribus populi et a senioribus sacerdotum 19:2

19:2 And go forth unto the valley of the son of Hinnom, which is by the entry of the east gate, and proclaim there the words that I shall tell thee,

et egredere ad uallem filii Ennom quae est iuxta introitum portae Fictilis et praedicabis ibi uerba quae ego loquar ad te 19:3 et dices audite uerbum Domini reges Iuda et habi- 19:3 And say, Hear ye the word of the LORD, O kings of

Judah, and inhabitants of Jerusalem; Thus saith the LORD of hosts, the God of Israel; Behold, I will bring evil upon this place, the which whosoever heareth, his ears shall tingle.

tatores Hierusalem haec dicit Dominus exercituum Deus Israhel ecce ego inducam adflictionem super locum istum ita ut omnis qui audierit illam tinniant aures eius 19:4 eo quod 19:4 Because they have forsaken me, and have estranged

this place, and have burned incense in it unto other gods, whom neither they nor their fathers have known, nor the kings of Judah, and have filled this place with the blood of innocents;

dereliquerint me et alienum fecerint locum istum et libauerint in eo diis alienis quos nescierunt ipsi et patres eorum et reges Iuda et repleuerunt locum istum sanguine innocentium 19:5 et aedificauerunt excelsa Baali ad conburendos 19:5 They have built also the high places of Baal, to burn

their sons with fire for burnt offerings unto Baal, which I commanded not, nor spake it, neither came it into my

filios suos igni in holocaustum Baali quae non praecepi nec mind: locutus sum nec ascenderunt in cor meum 19:6 propterea 19:6 Therefore, behold, the days come, saith the LORD,

that this place shall no more be called Tophet, nor The valley of the son of Hinnom, but The valley of slaughter.

ecce dies ueniunt dicit Dominus et non uocabitur locus iste amplius Thofeth et uallis filii Ennom sed uallis Occisionis 19:7 et dissipabo consilium Iudae et Hierusalem in loco isto 19:7 And I will make void the counsel of Judah and

Jerusalem in this place; and I will cause them to fall by the sword before their enemies, and by the hands of them that seek their lives: and their carcases will I give to be meat for the fowls of the heaven, and for the beasts of the earth.

et subuertam eos gladio in conspectu inimicorum suorum et in manu quaerentium animas eorum et dabo cadauera eorum escam uolatilibus caeli et bestiis terrae 19:8 et ponam ciui- 19:8 And I will make this city desolate, and an hissing;

every one that passeth thereby shall be astonished and hiss because of all the plagues thereof.

tatem hanc in stuporem et in sibilum omnis qui praeterierit per eam obstupescet et sibilabit super uniuersa plaga eius

Biblia Sacra Vulgata

19:9 And I will cause them to eat the flesh of their sons and 19:9 et cibabo eos carnibus filiorum suorum et carnibus fi-

the flesh of their daughters, and they shall eat every one the flesh of his friend in the siege and straitness, wherewith their enemies, and they that seek their lives, shall straiten them.

liarum suarum et unusquisque carnes amici sui comedet in obsidione et in angustia in qua concludent eos inimici eorum

19:10 Then shalt thou break the bottle in the sight of the et qui quaerunt animas eorum 19:10 et conteres lagunculam

men that go with thee,

19:11 And shalt say unto them, Thus saith the LORD of in oculis uirorum qui ibunt tecum 19:11 et dices ad eos haec

hosts; Even so will I break this people and this city, as one breaketh a potter’s vessel, that cannot be made whole again: and they shall bury them in Tophet, till there be no place to bury.

dicit Dominus exercituum sic conteram populum istum et ciuitatem istam sicut conteritur uas figuli quod non potest ultra instaurari et in Thofeth sepelientur eo quod non sit alius lo-

19:12 Thus will I do unto this place, saith the LORD, cus ad sepeliendum 19:12 sic faciam loco huic ait Dominus

and to the inhabitants thereof, and even make this city as Tophet:

et habitatoribus eius ut ponam ciuitatem istam sicut Thofeth

19:13 And the houses of Jerusalem, and the houses of the 19:13 et erunt domus Hierusalem et domus regum Iuda sicut

kings of Judah, shall be defiled as the place of Tophet, because of all the houses upon whose roofs they have burned incense unto all the host of heaven, and have poured out drink offerings unto other gods.

locus Thofeth inmundae omnes domus in quarum domatibus sacrificauerunt omni militiae caeli et libauerunt libamina diis

19:14 Then came Jeremiah from Tophet, whither the alienis 19:14 uenit autem Hieremias de Thofeth quo miserat

LORD had sent him to prophesy; and he stood in the court of the LORD’s house; and said to all the people,

eum Dominus ad prophetandum et stetit in atrio domus Do-

19:15 Thus saith the LORD of hosts, the God of Israel; mini et dixit ad omnem populum 19:15 haec dicit Dominus

Behold, I will bring upon this city and upon all her towns all the evil that I have pronounced against it, because they have hardened their necks, that they might not hear my words.

exercituum Deus Israhel ecce ego inducam super ciuitatem hanc et super omnes urbes eius uniuersa mala quae locutus sum aduersum eam quoniam indurauerunt ceruicem suam ut non audirent sermones meos

20:1 Now Pashur the son of Immer the priest, who was 20:1 et audiuit Phassur filius Emmer sacerdos qui constitu-

also chief governor in the house of the LORD, heard that Jeremiah prophesied these things.

tus erat princeps in domo Domini Hieremiam prophetantem

20:2 Then Pashur smote Jeremiah the prophet, and put him sermones istos 20:2 et percussit Phassur Hieremiam pro-

in the stocks that were in the high gate of Benjamin, which was by the house of the LORD.

phetam et misit eum in neruum quod erat in porta Beniamin

20:3 And it came to pass on the morrow, that Pashur superiori in domo Domini 20:3 cumque inluxisset in cras-

brought forth Jeremiah out of the stocks. Then said Jeremiah unto him, The LORD hath not called thy name

Pashur, but Magormissabib. tinum eduxit Phassur Hieremiam de neruo et dixit ad eum

Hieremias non Phassur uocauit Dominus nomen tuum sed

20:4 For thus saith the LORD, Behold, I will make thee a Pauorem undique 20:4 quia haec dicit Dominus ecce ego terror to thyself, and to all thy friends: and they shall fall by the sword of their enemies, and thine eyes shall behold it: and I will give all Judah into the hand of the king of Babylon, and he shall carry them captive into Babylon, and shall slay them with the sword.

dabo te in pauorem te et omnes amicos tuos et corruent gladio inimicorum suorum et oculi tui uidebunt et omnem Iudam dabo in manu regis Babylonis et traducet eos in Baby-

20:5 Moreover I will deliver all the strength of this city, lonem et percutiet eos gladio 20:5 et dabo uniuersam subsand all the labours thereof, and all the precious things thereof, and all the treasures of the kings of Judah will I give into the hand of their enemies, which shall spoil them, and take them, and carry them to Babylon.

tantiam ciuitatis huius et omnem laborem eius omneque pre-

Prophetia Ieremiae

tium et cunctos thesauros regum Iuda dabo in manu inimicorum eorum et diripient eos et tollent et ducent in Babylonem 20:6 tu autem Phassur et omnes habitatores domus tuae ibitis 20:6 And thou, Pashur, and all that dwell in thine house

shall go into captivity: and thou shalt come to Babylon, and there thou shalt die, and shalt be buried there, thou,

in captiuitatem et in Babylonem uenies et ibi morieris ibi- and all thy friends, to whom thou hast prophesied lies. que sepelieris tu et omnes amici tui quibus prophetasti mendacium 20:7 seduxisti me Domine et seductus sum fortior 20:7 O LORD, thou hast deceived me, and I was deceived;

thou art stronger than I, and hast prevailed: I am in derision daily, every one mocketh me.

me fuisti et inualuisti factus sum in derisum tota die omnes subsannant me 20:8 quia iam olim loquor uociferans 20:8 For since I spake, I cried out, I cried violence and

spoil; because the word of the LORD was made a reproach unto me, and a derision, daily.

iniquitatem et uastitatem clamito et factus est mihi sermo Domini in obprobrium et in derisum tota die 20:9 et dixi 20:9 Then I said, I will not make mention of him, nor

speak any more in his name. But his word was in mine heart as a burning fire shut up in my bones, and I was

non recordabor eius neque loquar ultra in nomine illius et weary with forbearing, and I could not stay. factus est in corde meo quasi ignis exaestuans claususque in ossibus meis et defeci ferre non sustinens 20:10 audiui 20:10 For I heard the defaming of many, fear on every side.

Report, say they, and we will report it. All my familiars watched for my halting, saying, Peradventure he will be enticed, and we shall prevail against him, and we shall take our revenge on him.

enim contumelias multorum et terrorem in circuitu persequimini et persequamur eum ab omnibus uiris qui erant pacifici mei et custodientes latus meum si quo modo decipiatur et praeualeamus aduersus eum et consequamur ultionem ex eo 20:11 Dominus autem mecum est quasi bellator fortis idcirco 20:11 But the LORD is with me as a mighty terrible one:

therefore my persecutors shall stumble, and they shall not prevail: they shall be greatly ashamed; for they shall not prosper: their everlasting confusion shall never be forgotten.

qui persequuntur me cadent et infirmi erunt confundentur uehementer quia non intellexerunt obprobrium sempiternum

quod numquam delebitur 20:12 et tu Domine exercituum 20:12 But, O LORD of hosts, that triest the righteous, and seest the reins and the heart, let me see thy vengeance on them: for unto thee have I opened my cause.

probator iusti qui uides renes et cor uideam quaeso ultionem tuam ex eis tibi enim reuelaui causam meam 20:13 cantate 20:13 Sing unto the LORD, praise ye the LORD: for he

hath delivered the soul of the poor from the hand of evildoers.

Domino laudate Dominum quia liberauit animam pauperis de manu malorum 20:14 maledicta dies in qua natus sum 20:14 Cursed be the day wherein I was born: let not the

day wherein my mother bare me be blessed.

dies in qua peperit me mater mea non sit benedicta 20:15 20:15 Cursed be the man who brought tidings to my father, saying, A man child is born unto thee; making him very glad.

maledictus uir qui adnuntiauit patri meo dicens natus est tibi puer masculus et quasi gaudio laetificauit eum 20:16 sit 20:16 And let that man be as the cities which the LORD

overthrew, and repented not: and let him hear the cry in the morning, and the shouting at noontide;

homo ille ut sunt ciuitates quas subuertit Dominus et non paenituit eum audiat clamorem mane et ululatum in tempore meridiano 20:17 qui non me interfecit a uulua ut fieret 20:17 Because he slew me not from the womb; or that my

mother might have been my grave, and her womb to be always great with me.

mihi mater mea sepulchrum et uulua eius conceptus aeternus 20:18 quare de uulua egressus sum ut uiderem laborem 20:18 Wherefore came I forth out of the womb to see

labour and sorrow, that my days should be consumed with shame?

Biblia Sacra Vulgata

et dolorem et consumerentur in confusione dies mei

21:1 The word which came unto Jeremiah from the 21:1 uerbum quod factum est ad Hieremiam a Domino

LORD, when king Zedekiah sent unto him Pashur the son of Melchiah, and Zephaniah the son of Maaseiah the

priest, saying, quando misit ad eum rex Sedecias Phassur filium Melchiae 21:2 Enquire, I pray thee, of the LORD for us; for Neb- et Sophoniam filium Maasiae sacerdotem dicens 21:2 inter-

uchadrezzar king of Babylon maketh war against us; if so be that the LORD will deal with us according to all his

wondrous works, that he may go up from us. roga pro nobis Dominum quia Nabuchodonosor rex Babylo-

nis proeliatur aduersum nos si forte faciat Dominus nobis-

21:3 Then said Jeremiah unto them, Thus shall ye say to cum secundum omnia mirabilia sua et recedat a nobis 21:3

Zedekiah:

21:4 Thus saith the LORD God of Israel; Behold, I will et dixit Hieremias ad eos sic dicetis Sedeciae 21:4 haec

turn back the weapons of war that are in your hands, wherewith ye fight against the king of Babylon, and against the Chaldeans, which besiege you without the walls, and I will assemble them into the midst of this city.

dicit Dominus Deus Israhel ecce ego conuertam uasa belli quae in manibus uestris sunt et quibus uos pugnatis aduersum regem Babylonis et Chaldeos qui obsident uos in circuitu murorum et congregabo ea in medio ciuitatis huius

21:5 And I myself will fight against you with an out- 21:5 et debellabo ego uos in manu extenta et brachio forti

stretched hand and with a strong arm, even in anger, and in fury, and in great wrath.

et in furore et in indignatione et in ira grandi 21:6 et per-

21:6 And I will smite the inhabitants of this city, both man and beast: they shall die of a great pestilence.

cutiam habitatores ciuitatis huius homines et bestiae pesti-

21:7 And afterward, saith the LORD, I will deliver lentia magna morientur 21:7 et post haec ait Dominus dabo

Zedekiah king of Judah, and his servants, and the people, and such as are left in this city from the pestilence, from the sword, and from the famine, into the hand of Nebuchadrezzar king of Babylon, and into the hand of their enemies, and into the hand of those that seek their life: and he shall smite them with the edge of the sword; he shall not spare them, neither have pity, nor have mercy.

Sedeciam regem Iuda et seruos eius et populum eius et qui derelicti sunt in ciuitate hac a peste et gladio et fame in manu Nabuchodonosor regis Babylonis et in manu inimicorum eorum et in manu quaerentium animam eorum et percutiet eos in ore gladii et non mouebitur neque parcet nec miserebitur

21:8 And unto this people thou shalt say, Thus saith the 21:8 et ad populum hunc dices haec dicit Dominus ecce ego

LORD; Behold, I set before you the way of life, and the way of death.

do coram uobis uiam uitae et uiam mortis 21:9 qui habita-

21:9 He that abideth in this city shall die by the sword, and by the famine, and by the pestilence: but he that goeth out, and falleth to the Chaldeans that besiege you, he shall live, and his life shall be unto him for a prey.

uerit in urbe hac morietur gladio et fame et peste qui autem egressus fuerit et transfugerit ad Chaldeos qui obsident uos

21:10 For I have set my face against this city for evil, and uiuet et erit ei anima sua quasi spolium 21:10 posui enim not for good, saith the LORD: it shall be given into the hand of the king of Babylon, and he shall burn it with fire.

faciem meam super ciuitatem hanc in malum et non in bonum ait Dominus in manu regis Babylonis dabitur et exuret

21:11 And touching the house of the king of Judah, say, eam igni 21:11 et domui regis Iuda audite uerbum Domini

Hear ye the word of the LORD;

21:12 O house of David, thus saith the LORD; Execute 21:12 domus Dauid haec dicit Dominus iudicate mane iudi-

judgment in the morning, and deliver him that is spoiled out of the hand of the oppressor, lest my fury go out like fire, and burn that none can quench it, because of the evil of your doings.

cium et eruite ui oppressum de manu calumniantis ne forte egrediatur ut ignis indignatio mea et succendatur et non sit

21:13 Behold, I am against thee, O inhabitant of the valley, qui extinguat propter malitiam studiorum uestrorum 21:13 and rock of the plain, saith the LORD; which say, Who shall come down against us? or who shall enter into our habitations?

Prophetia Ieremiae

ecce ego ad te habitatricem uallis solidae atque campestris ait Dominus qui dicitis quis percutiet nos et quis ingredietur domos nostras 21:14 et uisitabo super uos iuxta fructum 21:14 But I will punish you according to the fruit of your

doings, saith the LORD: and I will kindle a fire in the forest thereof, and it shall devour all things round about it.

studiorum uestrorum dicit Dominus et succendam ignem in saltu eius et deuorabit omnia in circuitu eius

22:1 haec dicit Dominus descende in domum regis Iuda et 22:1 Thus saith the LORD; Go down to the house of the

king of Judah, and speak there this word,

loqueris ibi uerbum hoc 22:2 et dices audi uerbum Domini 22:2 And say, Hear the word of the LORD, O king of Ju-

dah, that sittest upon the throne of David, thou, and thy servants, and thy people that enter in by these gates:

rex Iuda qui sedes super solium Dauid tu et serui tui et populus tuus qui ingredimini per portas istas 22:3 haec dicit Do- 22:3 Thus saith the LORD; Execute ye judgment and

righteousness, and deliver the spoiled out of the hand of the oppressor: and do no wrong, do no violence to the stranger, the fatherless, nor the widow, neither shed innocent blood in this place.

minus facite iudicium et iustitiam et liberate ui oppressum de manu calumniatoris et aduenam et pupillum et uiduam nolite contristare neque opprimatis inique et sanguinem innocentem ne effundatis in loco isto 22:4 si enim facientes feceri- 22:4 For if ye do this thing indeed, then shall there enter

in by the gates of this house kings sitting upon the throne of David, riding in chariots and on horses, he, and his ser-

tis uerbum istud ingredientur per portas domus huius reges vants, and his people. sedentes de genere Dauid super thronum eius et ascendentes currus et equos ipsi et serui et populus eorum 22:5 quod 22:5 But if ye will not hear these words, I swear by myself,

saith the LORD, that this house shall become a desolation.

si non audieritis uerba haec in memet ipso iuraui dicit Dominus quia in solitudinem erit domus haec 22:6 quia haec 22:6 For thus saith the LORD unto the king’s house of

Judah; Thou art Gilead unto me, and the head of Lebanon: yet surely I will make thee a wilderness, and cities which

dicit Dominus super domum regis Iuda Galaad tu mihi ca- are not inhabited. put Libani si non posuero te solitudinem urbes inhabitabiles 22:7 et sanctificabo super te interficientem uirum et arma eius 22:7 And I will prepare destroyers against thee, every one

with his weapons: and they shall cut down thy choice cedars, and cast them into the fire.

et succident electam cedrum tuam et praecipitabunt in ignem 22:8 et pertransibunt gentes multae per ciuitatem hanc 22:8 And many nations shall pass by this city, and they

shall say every man to his neighbour, Wherefore hath the LORD done thus unto this great city?

et dicet unusquisque proximo suo quare fecit Dominus sic ciuitati huic grandi 22:9 et respondebunt eo quod dereli- 22:9 Then they shall answer, Because they have forsaken

the covenant of the LORD their God, and worshipped other gods, and served them.

querint pactum Domini Dei sui et adorauerint deos alienos et seruierint eis 22:10 nolite flere mortuum neque lugeatis 22:10 Weep ye not for the dead, neither bemoan him: but

weep sore for him that goeth away: for he shall return no more, nor see his native country.

super eum fletu plangite eum qui egreditur quia non reuertetur ultra nec uidebit terram natiuitatis suae 22:11 quia haec 22:11 For thus saith the LORD touching Shallum the son

of Josiah king of Judah, which reigned instead of Josiah his father, which went forth out of this place; He shall not

dicit Dominus ad Sellum filium Iosiae regem Iuda qui reg- return thither any more: nauit pro Iosia patre suo qui egressus est de loco isto non reuertetur huc amplius 22:12 sed in loco ad quem transtuli 22:12 But he shall die in the place whither they have led

him captive, and shall see this land no more.

eum ibi morietur et terram istam non uidebit amplius 22:13 22:13 Woe unto him that buildeth his house by unrigh-

teousness, and his chambers by wrong; that useth his neighbour’s service without wages, and giveth him not for his work;

Biblia Sacra Vulgata

uae qui aedificat domum suam in iniustitia et cenacula sua non in iudicio amicum suum opprimet frustra et mercedem

22:14 That saith, I will build me a wide house and large eius non reddet ei 22:14 qui dicit aedificabo mihi domum

chambers, and cutteth him out windows; and it is cieled with cedar, and painted with vermilion.

latam et cenacula spatiosa qui aperit sibi fenestras et facit

22:15 Shalt thou reign, because thou closest thyself in laquearia cedrina pingitque sinopide 22:15 numquid regna-

cedar? did not thy father eat and drink, and do judgment and justice, and then it was well with him?

bis quoniam confers te cedro pater tuus numquid non comedit et bibit et fecit iudicium et iustitiam tunc cum bene

22:16 He judged the cause of the poor and needy; then it erat ei 22:16 iudicauit causam pauperis et egeni in bonum

was well with him: was not this to know me? saith the LORD.

suum numquid non ideo quia cognouit me dicit Dominus

22:17 But thine eyes and thine heart are not but for thy 22:17 tui uero oculi et cor ad auaritiam et ad sanguinem in-

covetousness, and for to shed innocent blood, and for oppression, and for violence, to do it.

nocentem fundendum et ad calumniam et ad cursum mali

22:18 Therefore thus saith the LORD concerning Je- operis 22:18 propterea haec dicit Dominus ad Ioachim fi-

hoiakim the son of Josiah king of Judah; They shall not lament for him, saying, Ah my brother! or, Ah sister! they shall not lament for him, saying, Ah lord! or, Ah his glory! lium Iosiae regem Iuda non plangent eum uae frater et uae

22:19 He shall be buried with the burial of an ass, drawn fratres non concrepabunt ei uae domine et uae inclite 22:19

and cast forth beyond the gates of Jerusalem.

sepultura asini sepelietur putrefactus et proiectus extra por-

22:20 Go up to Lebanon, and cry; and lift up thy voice in tas Hierusalem 22:20 ascende Libanum et clama et in Ba-

Bashan, and cry from the passages: for all thy lovers are destroyed.

san da uocem tuam et clama ad transeuntes quia contriti sunt

22:21 I spake unto thee in thy prosperity; but thou saidst, omnes amatores tui 22:21 locutus sum ad te in abundantia I will not hear. This hath been thy manner from thy youth, that thou obeyedst not my voice.

tua dixisti non audiam haec est uia tua ab adulescentia tua

22:22 The wind shall eat up all thy pastors, and thy lovers quia non audisti uocem meam 22:22 omnes pastores tuos shall go into captivity: surely then shalt thou be ashamed and confounded for all thy wickedness.

pascet uentus et amatores tui in captiuitatem ibunt et tunc

22:23 O inhabitant of Lebanon, that makest thy nest in the confunderis et erubesces ab omni malitia tua 22:23 quae se-

cedars, how gracious shalt thou be when pangs come upon thee, the pain as of a woman in travail!

des in Libano et nidificas in cedris quomodo congemuisti

22:24 As I live, saith the LORD, though Coniah the son cum uenissent tibi dolores quasi dolores parturientis 22:24

of Jehoiakim king of Judah were the signet upon my right hand, yet would I pluck thee thence;

uiuo ego dicit Dominus quia si fuerit Iechonias filius Ioachim regis Iuda anulus in manu dextera mea inde auellam

22:25 And I will give thee into the hand of them that seek eum 22:25 et dabo te in manu quaerentium animam tuam thy life, and into the hand of them whose face thou fearest, even into the hand of Nebuchadrezzar king of Babylon,

and into the hand of the Chaldeans. et in manu quorum tu formidas faciem et in manu Nabucho-22:26 And I will cast thee out, and thy mother that bare donosor regis Babylonis et in manu Chaldeorum 22:26 et

thee, into another country, where ye were not born; and there shall ye die.

mittam te et matrem tuam quae genuit te in terram alienam

22:27 But to the land whereunto they desire to return, in qua nati non estis ibique moriemini 22:27 et in terram ad

thither shall they not return.

quam ipsi leuant animam suam ut reuertantur illuc non re-

22:28 Is this man Coniah a despised broken idol? is he a uertentur 22:28 numquid uas fictile atque contritum uir iste

vessel wherein is no pleasure? wherefore are they cast out, he and his seed, and are cast into a land which they know not?

Prophetia Ieremiae

Iechonias numquid uas absque omni uoluptate quare abiecti sunt ipse et semen eius et proiecti in terram quam ignorauerunt 22:29 terra terra terra audi sermonem Domini 22:30 22:29 O earth, earth, earth, hear the word of the LORD.

22:30 Thus saith the LORD, Write ye this man childless, a man that shall not prosper in his days: for no man of his seed shall prosper, sitting upon the throne of David, and ruling any more in Judah.

haec dicit Dominus scribe uirum istum sterilem uirum qui in diebus suis non prosperabitur nec enim erit de semine eius uir qui sedeat super solium Dauid et potestatem habeat ultra in Iuda

23:1 uae pastoribus qui disperdunt et dilacerant gregem 23:1 Woe be unto the pastors that destroy and scatter the

sheep of my pasture! saith the LORD.

pascuae meae dicit Dominus 23:2 ideo haec dicit Dominus 23:2 Therefore thus saith the LORD God of Israel against

the pastors that feed my people; Ye have scattered my flock, and driven them away, and have not visited them: behold, I will visit upon you the evil of your doings, saith the LORD.

Deus Israhel ad pastores qui pascunt populum meum uos dispersistis gregem meum eiecistis eos et non uisitastis eos ecce ego uisitabo super uos malitiam studiorum uestrorum ait Dominus 23:3 et ego congregabo reliquias gregis mei 23:3 And I will gather the remnant of my flock out of all

countries whither I have driven them, and will bring them again to their folds; and they shall be fruitful and increase.

de omnibus terris ad quas eiecero eos illuc et conuertam eos ad rura sua et crescent et multiplicabuntur 23:4 et susci- 23:4 And I will set up shepherds over them which shall

feed them: and they shall fear no more, nor be dismayed, neither shall they be lacking, saith the LORD.

tabo super eos pastores et pascent eos non formidabunt ultra et non pauebunt et nullus quaeretur ex numero dicit Dominus 23:5 ecce dies ueniunt ait Dominus et suscitabo Dauid 23:5 Behold, the days come, saith the LORD, that I will

raise unto David a righteous Branch, and a King shall reign and prosper, and shall execute judgment and justice

germen iustum et regnabit rex et sapiens erit et faciet iu- in the earth. dicium et iustitiam in terra 23:6 in diebus illius saluabitur 23:6 In his days Judah shall be saved, and Israel shall

dwell safely: and this is his name whereby he shall be called, THE LORD OUR RIGHTEOUSNESS.

Iuda et Israhel habitabit confidenter et hoc est nomen quod uocabunt eum Dominus iustus noster 23:7 propter hoc ecce 23:7 Therefore, behold, the days come, saith the LORD,

that they shall no more say, The LORD liveth, which brought up the children of Israel out of the land of Egypt;

dies ueniunt dicit Dominus et non dicent ultra uiuit Dominus qui eduxit filios Israhel de terra Aegypti 23:8 sed uiuit 23:8 But, The LORD liveth, which brought up and which

led the seed of the house of Israel out of the north country, and from all countries whither I had driven them; and they

Dominus qui eduxit et adduxit semen domus Israhel de terra shall dwell in their own land. aquilonis et de cunctis terris ad quas eieceram eos illuc et habitabunt in terra sua 23:9 ad prophetas contritum est cor 23:9 Mine heart within me is broken because of the

prophets; all my bones shake; I am like a drunken man, and like a man whom wine hath overcome, because of the

meum in medio mei contremuerunt omnia ossa mea factus LORD, and because of the words of his holiness. sum quasi uir ebrius et quasi homo madidus a uino a facie Domini et a facie uerborum sanctorum eius 23:10 quia ad- 23:10 For the land is full of adulterers; for because of

swearing the land mourneth; the pleasant places of the wilderness are dried up, and their course is evil, and their

ulteris repleta est terra quia a facie maledictionis luxit terra force is not right. arefacta sunt arua deserti factus est cursus eorum malus et fortitudo eorum dissimilis 23:11 propheta namque et sacer- 23:11 For both prophet and priest are profane; yea, in my

house have I found their wickedness, saith the LORD.

Biblia Sacra Vulgata

dos polluti sunt et in domo mea inueni malum eorum ait Do-

23:12 Wherefore their way shall be unto them as slippery minus 23:12 idcirco uia eorum erit quasi lubricum in tene-

ways in the darkness: they shall be driven on, and fall therein: for I will bring evil upon them, even the year of

their visitation, saith the LORD. bris inpellentur enim et corruent in ea adferam enim super 23:13 And I have seen folly in the prophets of Samaria; eos mala annum uisitationis eorum ait Dominus 23:13 et

they prophesied in Baal, and caused my people Israel to err.

in prophetis Samariae uidi fatuitatem prophetabant in Baal

23:14 I have seen also in the prophets of Jerusalem an hor- et decipiebant populum meum Israhel 23:14 et in prophetis

rible thing: they commit adultery, and walk in lies: they strengthen also the hands of evildoers, that none doth return from his wickedness; they are all of them unto me as Sodom, and the inhabitants thereof as Gomorrah.

Hierusalem uidi similitudinem adulterium et iter mendacii et confortauerunt manus pessimorum ut non conuerteretur unusquisque a malitia sua facti sunt mihi omnes Sodoma et

23:15 Therefore thus saith the LORD of hosts concern- habitatores eius quasi Gomorra 23:15 propterea haec dicit

ing the prophets; Behold, I will feed them with wormwood, and make them drink the water of gall: for from the prophets of Jerusalem is profaneness gone forth into all the land.

Dominus exercituum ad prophetas ecce ego cibabo eos absinthio et potabo eos felle a prophetis enim Hierusalem est

23:16 Thus saith the LORD of hosts, Hearken not unto the egressa pollutio super omnem terram 23:16 haec dicit Do-

words of the prophets that prophesy unto you: they make you vain: they speak a vision of their own heart, and not

out of the mouth of the LORD. minus exercituum nolite audire uerba prophetarum qui pro-

phetant uobis et decipiunt uos uisionem cordis sui loquun-

23:17 They say still unto them that despise me, The LORD tur non de ore Domini 23:17 dicunt his qui blasphemant hath said, Ye shall have peace; and they say unto every one that walketh after the imagination of his own heart, No evil

shall come upon you. me locutus est Dominus pax erit uobis et omni qui ambu-

lat in prauitate cordis sui dixerunt non ueniet super uos ma-

23:18 For who hath stood in the counsel of the LORD, and lum 23:18 quis enim adfuit in consilio Domini et uidit et

hath perceived and heard his word? who hath marked his word, and heard it?

audiuit sermonem eius quis considerauit uerbum illius et au-

23:19 Behold, a whirlwind of the LORD is gone forth in diuit 23:19 ecce turbo dominicae indignationis egredietur

fury, even a grievous whirlwind: it shall fall grievously upon the head of the wicked.

et tempestas erumpens super caput impiorum ueniet 23:20

23:20 The anger of the LORD shall not return, until he have executed, and till he have performed the thoughts of his heart: in the latter days ye shall consider it perfectly.

non reuertetur furor Domini usque dum faciat et usque dum conpleat cogitationem cordis sui in nouissimis diebus intel-

23:21 I have not sent these prophets, yet they ran: I have legetis consilium eius 23:21 non mittebam prophetas et ipsi

not spoken to them, yet they prophesied.

23:22 But if they had stood in my counsel, and had caused currebant non loquebar ad eos et ipsi prophetabant 23:22 si

my people to hear my words, then they should have turned them from their evil way, and from the evil of their doings.

stetissent in consilio meo et nota fecissent uerba mea populo meo auertissem utique eos a uia sua mala et a pessimis co-

23:23 Am I a God at hand, saith the LORD, and not a God gitationibus suis 23:23 putasne Deus e uicino ego sum dicit

afar off? 23:24 Can any hide himself in secret places that I shall not Dominus et non Deus de longe 23:24 si occultabitur uir in see him? saith the LORD. Do not I fill heaven and earth? saith the LORD.

absconditis et ego non uidebo eum dicit Dominus numquid

23:25 I have heard what the prophets said, that prophesy non caelum et terram ego impleo ait Dominus 23:25 audiui lies in my name, saying, I have dreamed, I have dreamed.

quae dixerunt prophetae prophetantes in nomine meo men-

Prophetia Ieremiae

dacium atque dicentes somniaui somniaui 23:26 usquequo 23:26 How long shall this be in the heart of the prophets

that prophesy lies? yea, they are prophets of the deceit of their own heart;

istud in corde est prophetarum uaticinantium mendacium et prophetantium seductiones cordis sui 23:27 qui uolunt fa- 23:27 Which think to cause my people to forget my name

by their dreams which they tell every man to his neighbour, as their fathers have forgotten my name for Baal.

cere ut obliuiscatur populus meus nominis mei propter somnia eorum quae narrant unusquisque ad proximum suum sicut obliti sunt patres eorum nominis mei propter Baal 23:28 23:28 The prophet that hath a dream, let him tell a dream;

and he that hath my word, let him speak my word faithfully. What is the chaff to the wheat? saith the LORD.

propheta qui habet somnium narret somnium et qui habet sermonem meum loquatur sermonem meum uere quid paleis ad triticum dicit Dominus 23:29 numquid non uerba mea 23:29 Is not my word like as a fire? saith the LORD; and

like a hammer that breaketh the rock in pieces?

sunt quasi ignis ait Dominus et quasi malleus conterens pet-

ram 23:30 propterea ecce ego ad prophetas ait Dominus qui 23:30 Therefore, behold, I am against the prophets, saith the LORD, that steal my words every one from his neighbour.

furantur uerba mea unusquisque a proximo suo 23:31 ecce

23:31 Behold, I am against the prophets, saith the LORD, that use their tongues, and say, He saith.

ego ad prophetas ait Dominus qui adsumunt linguas suas et aiunt dicit Dominus 23:32 ecce ego ad prophetas somnian- 23:32 Behold, I am against them that prophesy false

dreams, saith the LORD, and do tell them, and cause my people to err by their lies, and by their lightness; yet I sent them not, nor commanded them: therefore they shall not profit this people at all, saith the LORD.

tes mendacium ait Dominus qui narrauerunt ea et seduxerunt populum meum in mendacio suo et in miraculis suis cum ego non misissem eos nec mandassem eis qui nihil profuerunt populo huic dicit Dominus 23:33 si igitur interrogauerit te 23:33 And when this people, or the prophet, or a priest,

shall ask thee, saying, What is the burden of the LORD? thou shalt then say unto them, What burden? I will even

populus iste uel propheta aut sacerdos dicens quod est onus forsake you, saith the LORD. Domini dices ad eos ut quid uobis onus proiciam quippe uos dicit Dominus 23:34 et prophetes et sacerdos et populus qui 23:34 And as for the prophet, and the priest, and the peo-

ple, that shall say, The burden of the LORD, I will even punish that man and his house.

dicit onus Domini uisitabo super uirum illum et super domum eius 23:35 haec dicetis unusquisque ad proximum et 23:35 Thus shall ye say every one to his neighbour, and

every one to his brother, What hath the LORD answered? and, What hath the LORD spoken?

ad fratrem suum quid respondit Dominus et quid locutus est Dominus 23:36 et onus Domini ultra non memorabitur quia 23:36 And the burden of the LORD shall ye mention no

more: for every man’s word shall be his burden; for ye have perverted the words of the living God, of the LORD

onus erit unicuique sermo suus et peruertitis uerba Dei ui- of hosts our God. uentis Domini exercituum Dei nostri 23:37 haec dices ad 23:37 Thus shalt thou say to the prophet, What hath the

LORD answered thee? and, What hath the LORD spoken?

prophetam quid respondit tibi Dominus et quid locutus est Dominus 23:38 si autem onus Domini dixeritis propter hoc 23:38 But since ye say, The burden of the LORD; there-

fore thus saith the LORD; Because ye say this word, The burden of the LORD, and I have sent unto you, saying, Ye

haec dicit Dominus quia dixistis sermonem istum onus Do- shall not say, The burden of the LORD;

mini et misi ad uos dicens nolite dicere onus Domini 23:39 23:39 Therefore, behold, I, even I, will utterly forget you, and I will forsake you, and the city that I gave you and your fathers, and cast you out of my presence:

propterea ecce ego tollam uos portans et derelinquam uos et ciuitatem quam dedi uobis et patribus uestris a facie mea

Biblia Sacra Vulgata

23:40 And I will bring an everlasting reproach upon you, 23:40 et dabo uos in obprobrium sempiternum et in ignomi-

and a perpetual shame, which shall not be forgotten.

niam aeternam quae numquam obliuione delebitur

24:1 The LORD shewed me, and, behold, two baskets of 24:1 ostendit mihi Dominus et ecce duo calathi pleni fi-

figs were set before the temple of the LORD, after that Nebuchadrezzar king of Babylon had carried away captive Jeconiah the son of Jehoiakim king of Judah, and the princes of Judah, with the carpenters and smiths, from Jerusalem, and had brought them to Babylon.

cis positi ante templum Domini postquam transtulit Nabuchodonosor rex Babylonis Iechoniam filium Ioachim regem Iuda et principes eius et fabrum et inclusorem de Hierusalem

24:2 One basket had very good figs, even like the figs that et adduxit eos in Babylonem 24:2 calathus unus ficus bonas

are first ripe: and the other basket had very naughty figs, which could not be eaten, they were so bad.

habebat nimis ut solent ficus esse primi temporis et calathus unus ficus habebat malas nimis quae comedi non poterant

24:3 Then said the LORD unto me, What seest thou, eo quod essent malae 24:3 et dixit Dominus ad me quid tu

Jeremiah? And I said, Figs; the good figs, very good; and the evil, very evil, that cannot be eaten, they are so evil.

uides Hieremia et dixi ficus ficus bonas bonas ualde et malas malas ualde quae comedi non possunt eo quod sint malae

24:4 Again the word of the LORD came unto me, saying, 24:4 et factum est uerbum Domini ad me dicens 24:5 haec

24:5 Thus saith the LORD, the God of Israel; Like these good figs, so will I acknowledge them that are carried away captive of Judah, whom I have sent out of this place into the land of the Chaldeans for their good.

dicit Dominus Deus Israhel sicut ficus hae bonae sic cognoscam transmigrationem Iuda quam emisi de loco isto in

24:6 For I will set mine eyes upon them for good, and I terram Chaldeorum in bonum 24:6 et ponam oculos meos

will bring them again to this land: and I will build them, and not pull them down; and I will plant them, and not

pluck them up. super eos ad placandum et reducam eos in terram hanc et ae-

dificabo eos et non destruam et plantabo eos et non euellam

24:7 And I will give them an heart to know me, that I am 24:7 et dabo eis cor ut sciant me quia ego sum Dominus et

the LORD: and they shall be my people, and I will be their God: for they shall return unto me with their whole heart.

erunt mihi in populum et ego ero eis in Deum quia reuerten-

24:8 And as the evil figs, which cannot be eaten, they tur ad me in toto corde suo 24:8 et sicut ficus pessimae quae

are so evil; surely thus saith the LORD, So will I give Zedekiah the king of Judah, and his princes, and the residue of Jerusalem, that remain in this land, and them that dwell in the land of Egypt:

comedi non possunt eo quod sint malae haec dicit Dominus sic dabo Sedeciam regem Iuda et principes eius et reliquos de Hierusalem qui remanserunt in urbe hac et qui habitant in

24:9 And I will deliver them to be removed into all the terra Aegypti 24:9 et dabo eos in uexationem adflictionem-

kingdoms of the earth for their hurt, to be a reproach and a proverb, a taunt and a curse, in all places whither I shall

drive them. que omnibus regnis terrae in obprobrium et in parabolam et

in prouerbium et in maledictionem in uniuersis locis ad quos

24:10 And I will send the sword, the famine, and the pesti- eieci eos 24:10 et mittam in eis gladium et famem et pestem lence, among them, till they be consumed from off the land that I gave unto them and to their fathers.

donec consumantur de terra quam dedi eis et patribus eorum

25:1 The word that came to Jeremiah concerning all the 25:1 uerbum quod factum est ad Hieremiam de omni popeople of Judah in the fourth year of Jehoiakim the son of Josiah king of Judah, that was the first year of Neb-

uchadrezzar king of Babylon; pulo Iudae in anno quarto Ioachim filii Iosiae regis Iuda ipse 25:2 The which Jeremiah the prophet spake unto all the est annus primus Nabuchodonosor regis Babylonis 25:2

people of Judah, and to all the inhabitants of Jerusalem, saying,

quae locutus est Hieremias propheta ad omnem populum

Prophetia Ieremiae

Iuda et ad uniuersos habitatores Hierusalem dicens 25:3 25:3 From the thirteenth year of Josiah the son of Amon

king of Judah, even unto this day, that is the three and twentieth year, the word of the LORD hath come unto me, and I have spoken unto you, rising early and speaking; but ye have not hearkened.

a tertiodecimo anno Iosiae filii Amon regis Iuda usque ad diem hanc iste est tertius et uicesimus annus factum est uerbum Domini ad me et locutus sum ad uos de nocte consurgens et loquens et non audistis 25:4 et misit Dominus ad 25:4 And the LORD hath sent unto you all his servants the

prophets, rising early and sending them; but ye have not hearkened, nor inclined your ear to hear.

uos omnes seruos suos prophetas consurgens diluculo mittensque et non audistis neque inclinastis aures uestras ut audiretis 25:5 cum diceret reuertimini unusquisque a uia sua 25:5 They said, Turn ye again now every one from his

evil way, and from the evil of your doings, and dwell in the land that the LORD hath given unto you and to your

mala et a pessimis cogitationibus uestris et habitabitis in ter- fathers for ever and ever: ram quam dedit Dominus uobis et patribus uestris a saeculo et usque in saeculum 25:6 et nolite ire post deos alienos ut 25:6 And go not after other gods to serve them, and to

worship them, and provoke me not to anger with the works of your hands; and I will do you no hurt.

seruiatis eis adoretisque eos neque me ad iracundiam prouocetis in operibus manuum uestrarum et non adfligam uos 25:7 et non audistis me dicit Dominus ut me ad iracundiam 25:7 Yet ye have not hearkened unto me, saith the LORD;

that ye might provoke me to anger with the works of your hands to your own hurt.

prouocaretis in operibus manuum uestrarum in malum uestrum 25:8 propterea haec dicit Dominus exercituum pro 25:8 Therefore thus saith the LORD of hosts; Because ye

have not heard my words,

eo quod non audistis uerba mea 25:9 ecce ego mittam et 25:9 Behold, I will send and take all the families of the north, saith the LORD, and Nebuchadrezzar the king of Babylon, my servant, and will bring them against this land, and against the inhabitants thereof, and against all these nations round about, and will utterly destroy them, and make them an astonishment, and an hissing, and perpetual desolations.

adsumam uniuersas cognationes aquilonis ait Dominus et ad Nabuchodonosor regem Babylonis seruum meum et adducam eos super terram istam et super habitatores eius et super omnes nationes quae in circuitu illius sunt et interficiam eos et ponam eos in stuporem et in sibilum et in solitudines sempiternas 25:10 perdamque ex eis uocem gaudii et uo- 25:10 Moreover I will take from them the voice of mirth,

and the voice of gladness, the voice of the bridegroom, and the voice of the bride, the sound of the millstones, and the

cem laetitiae uocem sponsae et uocem sponsi uocem molae light of the candle. et lumen lucernae 25:11 et erit uniuersa terra eius in soli- 25:11 And this whole land shall be a desolation, and an

astonishment; and these nations shall serve the king of Babylon seventy years.

tudinem et in stuporem et seruient omnes gentes istae regi Babylonis septuaginta annis 25:12 cumque impleti fuerint 25:12 And it shall come to pass, when seventy years are

accomplished, that I will punish the king of Babylon, and that nation, saith the LORD, for their iniquity, and the land

anni septuaginta uisitabo super regem Babylonis et super of the Chaldeans, and will make it perpetual desolations. gentem illam dicit Dominus iniquitatem eorum et super terram Chaldeorum et ponam illam in solitudines sempiternas 25:13 et adducam super terram illam omnia uerba mea quae 25:13 And I will bring upon that land all my words which

I have pronounced against it, even all that is written in this book, which Jeremiah hath prophesied against all the

locutus sum contra eam omne quod scriptum est in libro isto nations. quaecumque prophetauit Hieremias aduersum omnes gentes

Biblia Sacra Vulgata

25:14 For many nations and great kings shall serve them- 25:14 quia seruierunt eis cum essent gentes multae et reges

selves of them also: and I will recompense them according to their deeds, and according to the works of their own

hands. magni et reddam eis secundum opera eorum et secundum 25:15 For thus saith the LORD God of Israel unto me; facta manuum suarum 25:15 quia sic dicit Dominus exerci-

Take the wine cup of this fury at my hand, and cause all the nations, to whom I send thee, to drink it.

tuum Deus Israhel sume calicem uini furoris huius de manu mea et propinabis de illo cunctis gentibus ad quas ego mit-

25:16 And they shall drink, and be moved, and be mad, tam te 25:16 et bibent et turbabuntur et insanient a facie

because of the sword that I will send among them.

25:17 Then took I the cup at the LORD’s hand, and made gladii quem ego mittam inter eos 25:17 et accepi calicem

all the nations to drink, unto whom the LORD had sent me:

de manu Domini et propinaui cunctis gentibus ad quas misit

25:18 To wit, Jerusalem, and the cities of Judah, and the me Dominus 25:18 Hierusalem et ciuitatibus Iudae et regi-

kings thereof, and the princes thereof, to make them a desolation, an astonishment, an hissing, and a curse; as it is

this day; bus eius et principibus eius ut darem eos in solitudinem et in

stuporem in sibilum et in maledictionem sicut est dies ista

25:19 Pharaoh king of Egypt, and his servants, and his 25:19 Pharaoni regi Aegypti et seruis eius et principibus eius

princes, and all his people;

25:20 And all the mingled people, and all the kings of the et omni populo eius 25:20 et uniuersis generaliter cunctis

land of Uz, and all the kings of the land of the Philistines, and Ashkelon, and Azzah, and Ekron, and the remnant of

Ashdod, regibus terrae Ausitidis et cunctis regibus terrae Philisthim 25:21 Edom, and Moab, and the children of Ammon, et Ascaloni et Gazae et Accaroni et reliquiis Azoti 25:21

25:22 And all the kings of Tyrus, and all the kings of Zi- Idumeae et Moab et filiis Ammon 25:22 et cunctis regibus don, and the kings of the isles which are beyond the sea,

Tyri et cunctis regibus Sidonis et regibus terrae insularum

25:23 Dedan, and Tema, and Buz, and all that are in the qui sunt trans mare 25:23 et Dedan et Theman et Buz et utmost corners, 25:24 And all the kings of Arabia, and all the kings of the uniuersis qui adtonsi sunt in comam 25:24 et cunctis remingled people that dwell in the desert,

gibus Arabiae et cunctis regibus occidentis qui habitant in

25:25 And all the kings of Zimri, and all the kings of Elam, deserto 25:25 et cunctis regibus Zambri et cunctis regibus

and all the kings of the Medes, 25:26 And all the kings of the north, far and near, one with Aelam et cunctis regibus Medorum 25:26 et cunctis regianother, and all the kingdoms of the world, which are upon the face of the earth: and the king of Sheshach shall drink

after them. bus aquilonis de prope et de longe unicuique contra fratrem

suum et omnibus regnis terrae quae super faciem eius sunt

25:27 Therefore thou shalt say unto them, Thus saith the et rex Sesach bibet post eos 25:27 et dices ad eos haec di-LORD of hosts, the God of Israel; Drink ye, and be drunken, and spue, and fall, and rise no more, because of

the sword which I will send among you. cit Dominus exercituum Deus Israhel bibite et inebriamini

et uomite et cadite neque surgatis a facie gladii quem ego

25:28 And it shall be, if they refuse to take the cup at thine mittam inter uos 25:28 cumque noluerint accipere calicem hand to drink, then shalt thou say unto them, Thus saith the LORD of hosts; Ye shall certainly drink.

de manu ut bibant dices ad eos haec dicit Dominus exerci-

25:29 For, lo, I begin to bring evil on the city which is tuum bibentes bibetis 25:29 quia ecce in ciuitate in qua in-

called by my name, and should ye be utterly unpunished? Ye shall not be unpunished: for I will call for a sword upon all the inhabitants of the earth, saith the LORD of hosts. uocatum est nomen meum ego incipio adfligere et uos quasi

innocentes inmunes eritis non eritis inmunes gladium enim ego uoco super omnes habitatores terrae dicit Dominus ex-

Prophetia Ieremiae

ercituum 25:30 et tu prophetabis ad eos omnia uerba haec 25:30 Therefore prophesy thou against them all these

words, and say unto them, The LORD shall roar from on high, and utter his voice from his holy habitation; he shall mightily roar upon his habitation; he shall give a shout, as they that tread the grapes, against all the inhabitants of the earth.

et dices ad illos Dominus de excelso rugiet et de habitaculo sancto suo dabit uocem suam rugiens rugiet super decorem suum celeuma quasi calcantium concinetur aduersus omnes habitatores terrae 25:31 peruenit sonitus usque ad extrema 25:31 A noise shall come even to the ends of the earth;

for the LORD hath a controversy with the nations, he will plead with all flesh; he will give them that are wicked to

terrae quia iudicium Domino cum gentibus iudicatur ipse the sword, saith the LORD. cum omni carne impios tradidit gladio dicit Dominus 25:32 25:32 Thus saith the LORD of hosts, Behold, evil shall go

forth from nation to nation, and a great whirlwind shall be raised up from the coasts of the earth.

haec dicit Dominus exercituum ecce adflictio egredietur de gente in gentem et turbo magnus egredietur a summitatibus terrae 25:33 et erunt interfecti Domini in die illa a summo 25:33 And the slain of the LORD shall be at that day from

one end of the earth even unto the other end of the earth: they shall not be lamented, neither gathered, nor buried;

terrae usque ad summum eius non plangentur et non colli- they shall be dung upon the ground. gentur neque sepelientur in sterquilinium super faciem terrae iacebunt 25:34 ululate pastores et clamate et aspergite 25:34 Howl, ye shepherds, and cry; and wallow yourselves

in the ashes, ye principal of the flock: for the days of your slaughter and of your dispersions are accomplished; and

uos cinere optimates gregis quia conpleti sunt dies uestri ut ye shall fall like a pleasant vessel. interficiamini et dissipationes uestrae et cadetis quasi uasa pretiosa 25:35 et peribit fuga a pastoribus et saluatio ab op- 25:35 And the shepherds shall have no way to flee, nor the

principal of the flock to escape.

timatibus gregis 25:36 uox clamoris pastorum et ululatus 25:36 A voice of the cry of the shepherds, and an howling

of the principal of the flock, shall be heard: for the LORD hath spoiled their pasture.

optimatium gregis quia uastauit Dominus pascuam eorum 25:37 et conticuerunt arua pacis a facie irae furoris Domini 25:37 And the peaceable habitations are cut down because

of the fierce anger of the LORD.

25:38 dereliquit quasi leo umbraculum suum facta est terra 25:38 He hath forsaken his covert, as the lion: for their

land is desolate because of the fierceness of the oppressor, and because of his fierce anger.

eorum in desolationem a facie irae columbae et a facie irae furoris Domini

26:1 in principio regis Ioachim filii Iosiae regis Iuda fac- 26:1 In the beginning of the reign of Jehoiakim the son

of Josiah king of Judah came this word from the LORD, saying,

tum est uerbum istud a Domino dicens 26:2 haec dicit Do-

26:2 Thus saith the LORD; Stand in the court of the LORD’s house, and speak unto all the cities of Judah, which come to worship in the LORD’s house, all the words that I command thee to speak unto them; diminish not a word:

minus sta in atrio domus Domini et loqueris ad omnes ciuitates Iuda de quibus ueniunt ut adorent in domo Domini uniuersos sermones quos ego mandaui tibi ut loquaris ad eos noli subtrahere uerbum 26:3 si forte audiant et conuertantur 26:3 If so be they will hearken, and turn every man from

his evil way, that I may repent me of the evil, which I purpose to do unto them because of the evil of their doings.

unusquisque a uia sua mala et paeniteat me mali quod cogito facere eis propter malitias studiorum eorum 26:4 et dices ad 26:4 And thou shalt say unto them, Thus saith the LORD;

If ye will not hearken to me, to walk in my law, which I have set before you,

eos haec dicit Dominus si non audieritis me ut ambuletis in

lege mea quam dedi uobis 26:5 ut audiatis sermones ser- 26:5 To hearken to the words of my servants the prophets, whom I sent unto you, both rising up early, and sending them, but ye have not hearkened;

uorum meorum prophetarum quos ego misi ad uos de nocte

Biblia Sacra Vulgata

26:6 Then will I make this house like Shiloh, and will consurgens et dirigens et non audistis 26:6 dabo domum

make this city a curse to all the nations of the earth.

istam sicut Silo et urbem hanc dabo in maledictionem cunc-

26:7 So the priests and the prophets and all the people tis gentibus terrae 26:7 et audierunt sacerdotes et prophe-

heard Jeremiah speaking these words in the house of the LORD.

tae et omnis populus Hieremiam loquentem uerba haec in

26:8 Now it came to pass, when Jeremiah had made an domo Domini 26:8 cumque conplesset Hieremias loquens

end of speaking all that the LORD had commanded him to speak unto all the people, that the priests and the prophets

and all the people took him, saying, Thou shalt surely die. omnia quae praeceperat ei Dominus ut loqueretur ad uniuer-

sum populum adprehenderunt eum sacerdotes et prophetae

26:9 Why hast thou prophesied in the name of the LORD, et omnis populus dicens morte morietur 26:9 quare prophe-

saying, This house shall be like Shiloh, and this city shall be desolate without an inhabitant? And all the people were

gathered against Jeremiah in the house of the LORD. tauit in nomine Domini dicens sicut Silo erit domus haec et urbs ista desolabitur eo quod non sit habitator et congregatus est omnis populus aduersum Hieremiam in domum Domini

26:10 When the princes of Judah heard these things, then 26:10 et audierunt principes Iuda uerba haec et ascenderunt

they came up from the king’s house unto the house of the LORD, and sat down in the entry of the new gate of the

LORD’s house. de domo regis in domum Domini et sederunt in introitu por-26:11 Then spake the priests and the prophets unto the tae Domini nouae 26:11 et locuti sunt sacerdotes et prophe-

princes and to all the people, saying, This man is worthy to die; for he hath prophesied against this city, as ye have

heard with your ears. tae ad principes et ad omnem populum dicentes iudicium

mortis est uiro huic quia prophetauit aduersum ciuitatem is-

26:12 Then spake Jeremiah unto all the princes and to all tam sicut audistis auribus uestris 26:12 et ait Hieremias ad

the people, saying, The LORD sent me to prophesy against this house and against this city all the words that ye have

heard. omnes principes et ad uniuersum populum dicens Dominus

misit me ut prophetarem ad domum istam et ad ciuitatem

26:13 Therefore now amend your ways and your doings, hanc omnia uerba quae audistis 26:13 nunc ergo bonas fa-

and obey the voice of the LORD your God; and the LORD will repent him of the evil that he hath pronounced against

you. cite uias uestras et studia uestra et audite uocem Domini Dei

uestri et paenitebit Dominum mali quod locutus est aduer-

26:14 As for me, behold, I am in your hand: do with me sum uos 26:14 ego autem ecce in manibus uestris sum facite

as seemeth good and meet unto you.

26:15 But know ye for certain, that if ye put me to death, mihi ut bonum et rectum est in oculis uestris 26:15 uerum-

ye shall surely bring innocent blood upon yourselves, and upon this city, and upon the inhabitants thereof: for of a truth the LORD hath sent me unto you to speak all these words in your ears.

tamen scitote et cognoscite quod si occideritis me sanguinem innocentem traditis contra uosmet ipsos et contra ciuitatem istam et habitatores eius in ueritate enim misit me Dominus ad uos ut loquerer in auribus uestris omnia uerba haec

26:16 Then said the princes and all the people unto the 26:16 et dixerunt principes et omnis populus ad sacerdotes et

priests and to the prophets; This man is not worthy to die: for he hath spoken to us in the name of the LORD our

God. prophetas non est uiro huic iudicium mortis quia in nomine 26:17 Then rose up certain of the elders of the land, and Domini Dei nostri locutus est ad nos 26:17 surrexerunt ergo

spake to all the assembly of the people, saying,

uiri de senioribus terrae et dixerunt ad omnem coetum po-

26:18 Micah the Morasthite prophesied in the days of puli loquentes 26:18 Michas de Morasthim fuit propheta in

Hezekiah king of Judah, and spake to all the people of Judah, saying, Thus saith the LORD of hosts; Zion shall be plowed like a field, and Jerusalem shall become heaps, and the mountain of the house as the high places of a forest.

Prophetia Ieremiae

diebus Ezechiae regis Iudae et ait ad omnem populum Iudae dicens haec dicit Dominus exercituum Sion quasi ager arabitur et Hierusalem in aceruum lapidum erit et mons domus in excelsa siluarum 26:19 numquid morte condemnauit eum 26:19 Did Hezekiah king of Judah and all Judah put him

at all to death? did he not fear the LORD, and besought the LORD, and the LORD repented him of the evil which he had pronounced against them? Thus might we procure great evil against our souls.

Ezechias rex Iuda et omnis Iuda numquid non timuerunt Dominum et deprecati sunt faciem Domini et paenituit Dominum mali quod locutus erat aduersum eos itaque nos facimus malum grande contra animas nostras 26:20 fuit quoque uir 26:20 And there was also a man that prophesied in the

name of the LORD, Urijah the son of Shemaiah of Kirjathjearim, who prophesied against this city and against

prophetans in nomine Domini Vrias filius Semei de Cariat- this land according to all the words of Jeremiah. hiarim et prophetauit aduersum ciuitatem istam et aduersum terram hanc iuxta uniuersa uerba Hieremiae 26:21 et audi- 26:21 And when Jehoiakim the king, with all his mighty

men, and all the princes, heard his words, the king sought to put him to death: but when Urijah heard it, he was

uit rex Ioachim et omnes potentes et principes eius uerba afraid, and fled, and went into Egypt; haec et quaesiuit rex interficere eum et audiuit Vrias et timuit fugitque et ingressus est Aegyptum 26:22 et misit rex 26:22 And Jehoiakim the king sent men into Egypt,

namely, Elnathan the son of Achbor, and certain men with him into Egypt.

Ioachim uiros in Aegyptum Elnathan filium Achobor et uiros cum eo in Aegyptum 26:23 et eduxerunt Vriam de Ae- 26:23 And they fetched forth Urijah out of Egypt, and

brought him unto Jehoiakim the king; who slew him with the sword, and cast his dead body into the graves of the

gypto et adduxerunt eum ad regem Ioachim et percussit eum common people. gladio et proiecit cadauer eius in sepulchris uulgi ignobilis 26:24 igitur manus Ahicam filii Saphan fuit cum Hieremia ut 26:24 Nevertheless the hand of Ahikam the son of

Shaphan was with Jeremiah, that they should not give him into the hand of the people to put him to death.

non traderetur in manu populi et interficerent eum

27:1 in principio regni Ioachim filii Iosiae regis Iuda fac- 27:1 In the beginning of the reign of Jehoiakim the son of

Josiah king of Judah came this word unto Jeremiah from the LORD, saying,

tum est uerbum istud ad Hieremiam a Domino dicens 27:2

27:2 Thus saith the LORD to me; Make thee bonds and yokes, and put them upon thy neck,

haec dicit Dominus ad me fac tibi uincula et catenas et pones eas in collo tuo 27:3 et mittes eas ad regem Edom et ad 27:3 And send them to the king of Edom, and to the king

of Moab, and to the king of the Ammonites, and to the king of Tyrus, and to the king of Zidon, by the hand of the messengers which come to Jerusalem unto Zedekiah king of Judah;

regem Moab et ad regem filiorum Ammon et ad regem Tyri et ad regem Sidonis in manu nuntiorum qui uenerunt Hierusalem ad Sedeciam regem Iuda 27:4 et praecipies eis ut 27:4 And command them to say unto their masters, Thus

saith the LORD of hosts, the God of Israel; Thus shall ye say unto your masters;

ad dominos suos loquantur haec dicit Dominus exercituum Deus Israhel haec dicetis ad dominos uestros 27:5 ego feci 27:5 I have made the earth, the man and the beast that

are upon the ground, by my great power and by my outstretched arm, and have given it unto whom it seemed

terram et hominem et iumenta quae sunt super faciem terrae meet unto me. in fortitudine mea magna et in brachio meo extento et dedi eam ei qui placuit in oculis meis 27:6 et nunc itaque ego 27:6 And now have I given all these lands into the hand of

Nebuchadnezzar the king of Babylon, my servant; and the beasts of the field have I given him also to serve him.

dedi omnes terras istas in manu Nabuchodonosor regis Ba-

Biblia Sacra Vulgata

bylonis serui mei insuper et bestias agri dedi ei ut seruiant illi

27:7 And all nations shall serve him, and his son, and his 27:7 et seruient ei omnes gentes et filio eius et filio filii eius

son’s son, until the very time of his land come: and then many nations and great kings shall serve themselves of

him. donec ueniat tempus terrae eius et ipsius et seruient ei gentes

27:8 And it shall come to pass, that the nation and king- multae et reges magni 27:8 gens autem et regnum quod non dom which will not serve the same Nebuchadnezzar the king of Babylon, and that will not put their neck under the yoke of the king of Babylon, that nation will I punish, saith the LORD, with the sword, and with the famine, and with the pestilence, until I have consumed them by his hand.

seruierit Nabuchodonosor regi Babylonis et quicumque non curuauerit collum suum sub iugo regis Babylonis in gladio et in fame et in peste uisitabo super gentem illam ait Domi-

27:9 Therefore hearken not ye to your prophets, nor to nus donec consumam eos in manu eius 27:9 uos ergo nolite

your diviners, nor to your dreamers, nor to your enchanters, nor to your sorcerers, which speak unto you, say-

ing, Ye shall not serve the king of Babylon: audire prophetas uestros et diuinos et somniatores et augu-

res et maleficos qui dicunt uobis non seruietis regi Babylonis

27:10 For they prophesy a lie unto you, to remove you far 27:10 quia mendacium prophetant uobis ut longe faciant uos

from your land; and that I should drive you out, and ye should perish.

de terra uestra et eiciant uos et pereatis 27:11 porro gens

27:11 But the nations that bring their neck under the yoke of the king of Babylon, and serve him, those will I let remain still in their own land, saith the LORD; and they shall till it, and dwell therein.

quae subiecerit ceruicem suam sub iugo regis Babylonis et seruierit ei dimittam eam in terra sua dicit Dominus et colet

27:12 I spake also to Zedekiah king of Judah according to eam et habitabit in ea 27:12 et ad Sedeciam regem Iuda loall these words, saying, Bring your necks under the yoke of the king of Babylon, and serve him and his people, and

live. cutus sum secundum omnia uerba haec dicens subicite colla

uestra sub iugo regis Babylonis et seruite ei et populo eius

27:13 Why will ye die, thou and thy people, by the sword, et uiuetis 27:13 quare moriemini tu et populus tuus gladio

by the famine, and by the pestilence, as the LORD hath spoken against the nation that will not serve the king of

Babylon? fame et peste sicut locutus est Dominus ad gentem quae ser-27:14 Therefore hearken not unto the words of the uire noluerit regi Babylonis 27:14 nolite audire uerba pro-

prophets that speak unto you, saying, Ye shall not serve the king of Babylon: for they prophesy a lie unto you.

phetarum dicentium uobis non seruietis regi Babylonis quia

27:15 For I have not sent them, saith the LORD, yet they mendacium ipsi loquuntur uobis 27:15 quia non misi eos

prophesy a lie in my name; that I might drive you out, and that ye might perish, ye, and the prophets that prophesy

unto you. ait Dominus et ipsi prophetant in nomine meo mendaciter

ut eiciant uos et pereatis tam uos quam prophetae qui uati-

27:16 Also I spake to the priests and to all this people, cinantur uobis 27:16 et ad sacerdotes et ad populum istum

saying, Thus saith the LORD; Hearken not to the words of your prophets that prophesy unto you, saying, Behold, the vessels of the LORD’s house shall now shortly be brought again from Babylon: for they prophesy a lie unto you.

locutus sum dicens haec dicit Dominus nolite audire uerba prophetarum uestrorum qui prophetant uobis dicentes ecce uasa Domini reuertentur de Babylone nunc cito mendacium

27:17 Hearken not unto them; serve the king of Babylon, enim prophetant uobis 27:17 nolite ergo audire eos sed ser-

and live: wherefore should this city be laid waste?

uite regi Babylonis ut uiuatis quare datur haec ciuitas in soli-

27:18 But if they be prophets, and if the word of the tudinem 27:18 et si prophetae sunt et est uerbum Domini in

LORD be with them, let them now make intercession to the LORD of hosts, that the vessels which are left in the house of the LORD, and in the house of the king of Judah, and at Jerusalem, go not to Babylon.

eis occurrant Domino exercituum ut non ueniant uasa quae derelicta fuerant in domum Domini et in domum regis Iuda

Prophetia Ieremiae

et in Hierusalem in Babylonem 27:19 quia haec dicit Domi- 27:19 For thus saith the LORD of hosts concerning the

pillars, and concerning the sea, and concerning the bases, and concerning the residue of the vessels that remain in

nus exercituum ad columnas et ad mare et ad bases et ad re- this city. liqua uasorum quae remanserunt in ciuitate hac 27:20 quae 27:20 Which Nebuchadnezzar king of Babylon took not,

when he carried away captive Jeconiah the son of Jehoiakim king of Judah from Jerusalem to Babylon, and

non tulit Nabuchodonosor rex Babylonis cum transferret Ie- all the nobles of Judah and Jerusalem; choniam filium Ioachim regem Iuda de Hierusalem in Babylonem et omnes optimates Iuda et Hierusalem 27:21 quia 27:21 Yea, thus saith the LORD of hosts, the God of Is-

rael, concerning the vessels that remain in the house of the LORD, and in the house of the king of Judah and of

haec dicit Dominus exercituum Deus Israhel ad uasa quae Jerusalem; derelicta sunt in domum Domini et in domum regis Iuda et Hierusalem 27:22 in Babylonem transferentur et ibi erunt 27:22 They shall be carried to Babylon, and there shall

they be until the day that I visit them, saith the LORD; then will I bring them up, and restore them to this place.

usque ad diem uisitationis suae dicit Dominus et adferri faciam ea et restitui in loco isto

28:1 et factum est in anno illo in principio regni Sedeciae 28:1 And it came to pass the same year, in the beginning

of the reign of Zedekiah king of Judah, in the fourth year, and in the fifth month, that Hananiah the son of Azur the prophet, which was of Gibeon, spake unto me in the house of the LORD, in the presence of the priests and of all the people, saying,

regis Iuda in anno quarto in mense quinto dixit ad me Ananias filius Azur propheta de Gabaon in domo Domini coram sacerdotibus et omni populo dicens 28:2 haec dicit Domi- 28:2 Thus speaketh the LORD of hosts, the God of Israel,

saying, I have broken the yoke of the king of Babylon.

nus exercituum Deus Israhel contriui iugum regis Babylonis 28:3 adhuc duo anni dierum et ego referri faciam ad locum is- 28:3 Within two full years will I bring again into this place

all the vessels of the LORD’s house, that Nebuchadnezzar king of Babylon took away from this place, and carried

tum omnia uasa Domini quae tulit Nabuchodonosor rex Ba- them to Babylon: bylonis de loco isto et transtulit ea in Babylonem 28:4 et 28:4 And I will bring again to this place Jeconiah the son

of Jehoiakim king of Judah, with all the captives of Judah, that went into Babylon, saith the LORD: for I will break

Iechoniam filium Ioachim regem Iuda et omnem transmig- the yoke of the king of Babylon. rationem Iudae qui ingressi sunt in Babylonem ego conuertam ad locum istum ait Dominus conteram enim iugum regis Babylonis 28:5 et dixit Hieremias propheta ad Ananiam 28:5 Then the prophet Jeremiah said unto the prophet

Hananiah in the presence of the priests, and in the presence of all the people that stood in the house of the LORD,

prophetam in oculis sacerdotum et in oculis omnis populi qui stabant in domo Domini 28:6 et ait Hieremias propheta 28:6 Even the prophet Jeremiah said, Amen: the LORD do

so: the LORD perform thy words which thou hast prophesied, to bring again the vessels of the LORD’s house, and all that is carried away captive, from Babylon into this place.

amen sic faciat Dominus suscitet Dominus uerba tua quae prophetasti ut referantur uasa in domum Domini et omnis transmigratio de Babylone ad locum istum 28:7 uerumta- 28:7 Nevertheless hear thou now this word that I speak in

thine ears, and in the ears of all the people;

men audi uerbum hoc quod ego loquor in auribus tuis et in auribus uniuersi populi 28:8 prophetae qui fuerunt ante me 28:8 The prophets that have been before me and before

thee of old prophesied both against many countries, and against great kingdoms, of war, and of evil, and of pesti-

et te ab initio et prophetauerunt super terras multas et super lence. regna magna de proelio et de adflictione et de fame 28:9 28:9 The prophet which prophesieth of peace, when the

word of the prophet shall come to pass, then shall the prophet be known, that the LORD hath truly sent him.

propheta qui uaticinatus est pacem cum uenerit uerbum eius

Biblia Sacra Vulgata

28:10 Then Hananiah the prophet took the yoke from off scietur propheta quem misit Dominus in ueritate 28:10 et

the prophet Jeremiah’s neck, and brake it.

tulit Ananias propheta catenam de collo Hieremiae prophe-

28:11 And Hananiah spake in the presence of all the peo- tae et confregit eam 28:11 et ait Ananias in conspectu om-

ple, saying, Thus saith the LORD; Even so will I break the yoke of Nebuchadnezzar king of Babylon from the neck of all nations within the space of two full years. And the prophet Jeremiah went his way.

nis populi dicens haec dicit Dominus sic confringam iugum Nabuchodonosor regis Babylonis post duos annos dierum de

28:12 Then the word of the LORD came unto Jeremiah the collo omnium gentium 28:12 et abiit Hieremias prophetes

prophet, after that Hananiah the prophet had broken the yoke from off the neck of the prophet Jeremiah, saying,

in uiam suam et factum est uerbum Domini ad Hieremiam postquam confregit Ananias propheta catenam de collo Hie-

28:13 Go and tell Hananiah, saying, Thus saith the LORD; remiae prophetae dicens 28:13 uade et dices Ananiae haec

Thou hast broken the yokes of wood; but thou shalt make for them yokes of iron.

dicit Dominus catenas ligneas contriuisti et facies pro eis

28:14 For thus saith the LORD of hosts, the God of Israel; catenas ferreas 28:14 quia haec dicit Dominus exercituum

I have put a yoke of iron upon the neck of all these nations, that they may serve Nebuchadnezzar king of Babylon; and they shall serve him: and I have given him the beasts of the field also.

Deus Israhel iugum ferreum posui super collum cunctarum gentium istarum ut seruiant Nabuchodonosor regi Babylonis

28:15 Then said the prophet Jeremiah unto Hananiah the et seruient ei insuper et bestias terrae dedi ei 28:15 et dixit

prophet, Hear now, Hananiah; The LORD hath not sent thee; but thou makest this people to trust in a lie.

Hieremias propheta ad Ananiam prophetam audi Anania non misit te Dominus et tu confidere fecisti populum istum in

28:16 Therefore thus saith the LORD; Behold, I will cast mendacio 28:16 idcirco haec dicit Dominus ecce emittam thee from off the face of the earth: this year thou shalt die, because thou hast taught rebellion against the LORD.

te a facie terrae hoc anno morieris aduersum Dominum enim

28:17 So Hananiah the prophet died the same year in the locutus es 28:17 et mortuus est Ananias propheta in anno

seventh month.

illo mense septimo

29:1 Now these are the words of the letter that Jeremiah 29:1 et haec sunt uerba libri quae misit Hieremias prothe prophet sent from Jerusalem unto the residue of the elders which were carried away captives, and to the priests, and to the prophets, and to all the people whom Nebuchadnezzar had carried away captive from Jerusalem to Babylon;

pheta de Hierusalem ad reliquias seniorum transmigrationis et ad sacerdotes et ad prophetas et ad omnem populum quem transduxerat Nabuchodonosor de Hierusalem in Babylonem

29:2 (After that Jeconiah the king, and the queen, and the 29:2 postquam egressus est Iechonias rex et domina et eueunuchs, the princes of Judah and Jerusalem, and the carpenters, and the smiths, were departed from Jerusalem;)

nuchi et principes Iuda et Hierusalem et faber et inclusor de

29:3 By the hand of Elasah the son of Shaphan, and Hierusalem 29:3 in manu Ellasa filii Saphan et Gamaliae fi-

Gemariah the son of Hilkiah, (whom Zedekiah king of Judah sent unto Babylon to Nebuchadnezzar king of Baby-

lon) saying, lii Helciae quos misit Sedecias rex Iuda ad Nabuchodonosor

29:4 Thus saith the LORD of hosts, the God of Israel, unto regem Babylonis in Babylonem dicens 29:4 haec dicit Doall that are carried away captives, whom I have caused to be carried away from Jerusalem unto Babylon;

minus exercituum Deus Israhel omni transmigrationi quam

29:5 Build ye houses, and dwell in them; and plant gar- transtuli de Hierusalem in Babylonem 29:5 aedificate do-

dens, and eat the fruit of them;

mos et habitate et plantate hortos et comedite fructum eorum

29:6 Take ye wives, and beget sons and daughters; and 29:6 accipite uxores et generate filios et filias date filiis ue-

take wives for your sons, and give your daughters to husbands, that they may bear sons and daughters; that ye may be increased there, and not diminished.

Prophetia Ieremiae

stris uxores et filias uestras date uiris et pariant filios et filias et multiplicamini ibi et nolite esse pauci numero 29:7 29:7 And seek the peace of the city whither I have caused

you to be carried away captives, and pray unto the LORD for it: for in the peace thereof shall ye have peace.

et quaerite pacem ciuitatis ad quam transmigrare uos feci et orate pro ea ad Dominum quia in pace illius erit pax uobis 29:8 haec enim dicit Dominus exercituum Deus Israhel non 29:8 For thus saith the LORD of hosts, the God of Israel;

Let not your prophets and your diviners, that be in the midst of you, deceive you, neither hearken to your dreams

uos inducant prophetae uestri qui sunt in medio uestrum et which ye cause to be dreamed. diuini uestri et ne adtendatis ad somnia uestra quae uos somniatis 29:9 quia falso ipsi prophetant uobis in nomine meo 29:9 For they prophesy falsely unto you in my name: I

have not sent them, saith the LORD.

et non misi eos dicit Dominus 29:10 quia haec dicit Do- 29:10 For thus saith the LORD, That after seventy years be accomplished at Babylon I will visit you, and perform my good word toward you, in causing you to return to this

minus cum coeperint impleri in Babylone septuaginta anni place. uisitabo uos et suscitabo super uos uerbum meum bonum ut reducam uos ad locum istum 29:11 ego enim scio cogita- 29:11 For I know the thoughts that I think toward you,

saith the LORD, thoughts of peace, and not of evil, to give you an expected end.

tiones quas cogito super uos ait Dominus cogitationes pacis et non adflictionis ut dem uobis finem et patientiam 29:12 29:12 Then shall ye call upon me, and ye shall go and pray

unto me, and I will hearken unto you.

et inuocabitis me et ibitis et orabitis me et exaudiam uos 29:13 quaeretis me et inuenietis cum quaesieritis me in toto 29:13 And ye shall seek me, and find me, when ye shall

search for me with all your heart.

corde uestro 29:14 et inueniar a uobis ait Dominus et re- 29:14 And I will be found of you, saith the LORD: and I

will turn away your captivity, and I will gather you from all the nations, and from all the places whither I have driven you, saith the LORD; and I will bring you again into the place whence I caused you to be carried away captive.

ducam captiuitatem uestram et congregabo uos de uniuersis gentibus et de cunctis locis ad quae expuli uos dicit Dominus et reuerti uos faciam de loco ad quem transmigrare uos feci 29:15 quia dixistis suscitauit nobis Dominus prophetas in Ba- 29:15 Because ye have said, The LORD hath raised us up

prophets in Babylon;

bylone 29:16 quia haec dicit Dominus ad regem qui sedet 29:16 Know that thus saith the LORD of the king that sit-

teth upon the throne of David, and of all the people that dwelleth in this city, and of your brethren that are not gone

super solium Dauid et ad omnem populum habitatorem ur- forth with you into captivity; bis huius ad fratres uestros qui non sunt egressi uobiscum in transmigrationem 29:17 haec dicit Dominus exercituum 29:17 Thus saith the LORD of hosts; Behold, I will send

upon them the sword, the famine, and the pestilence, and will make them like vile figs, that cannot be eaten, they are

ecce mittam in eis gladium et famem et pestem et ponam so evil. eos quasi ficus malas quae comedi non possunt eo quod pessimae sint 29:18 et persequar eos in gladio in fame et in 29:18 And I will persecute them with the sword, with the

famine, and with the pestilence, and will deliver them to be removed to all the kingdoms of the earth, to be a curse, and an astonishment, and an hissing, and a reproach, among all the nations whither I have driven them:

pestilentia et dabo eos in uexationem uniuersis regnis terrae in maledictionem et in stuporem et in sibilum et in obprobrium cunctis gentibus ad quas ego eieci eos 29:19 eo 29:19 Because they have not hearkened to my words, saith

the LORD, which I sent unto them by my servants the prophets, rising up early and sending them; but ye would

quod non audierint uerba mea dicit Dominus quae misi ad not hear, saith the LORD. eos per seruos meos prophetas de nocte consurgens et mit-

Biblia Sacra Vulgata

29:20 Hear ye therefore the word of the LORD, all ye of tens et non audistis dicit Dominus 29:20 uos ergo audite

the captivity, whom I have sent from Jerusalem to Babylon:

uerbum Domini omnis transmigratio quam emisi de Hieru-

29:21 Thus saith the LORD of hosts, the God of Israel, of salem in Babylonem 29:21 haec dicit Dominus exercituum

Ahab the son of Kolaiah, and of Zedekiah the son of Maaseiah, which prophesy a lie unto you in my name; Behold, I will deliver them into the hand of Nebuchadrezzar king of Babylon; and he shall slay them before your eyes;

Deus Israhel ad Ahab filium Culia et ad Sedeciam filium Maasiae qui prophetant uobis in nomine meo mendaciter ecce ego tradam eos in manu Nabuchodonosor regis Babylonis

29:22 And of them shall be taken up a curse by all the cap- et percutiet eos in oculis uestris 29:22 et adsumetur ex eis

tivity of Judah which are in Babylon, saying, The LORD make thee like Zedekiah and like Ahab, whom the king of

Babylon roasted in the fire; maledictio omni transmigrationi Iuda quae est in Babylone

dicentium ponat te Dominus sicut Sedeciam et sicut Ahab

29:23 Because they have committed villany in Israel, and quos frixit rex Babylonis in igne 29:23 pro eo quod fecerint

have committed adultery with their neighbours’ wives, and have spoken lying words in my name, which I have not commanded them; even I know, and am a witness, saith the LORD.

stultitiam in Israhel et moechati sunt in uxores amicorum suorum et locuti sunt uerbum in nomine meo mendaciter quod non mandaui eis ego sum iudex et testis dicit Domi-

29:24 Thus shalt thou also speak to Shemaiah the Nehe- nus 29:24 et ad Semeiam Neelamiten dices 29:25 haec di-

lamite, saying, 29:25 Thus speaketh the LORD of hosts, the God of Israel, saying, Because thou hast sent letters in thy name unto all the people that are at Jerusalem, and to Zephaniah the son of Maaseiah the priest, and to all the priests, saying,

cit Dominus exercituum Deus Israhel pro eo quod misisti in nomine tuo libros ad omnem populum qui est in Hierusalem et ad Sophoniam filium Maasiae sacerdotem et ad uniuersos

29:26 The LORD hath made thee priest in the stead of Je- sacerdotes dicens 29:26 Dominus dedit te sacerdotem pro

hoiada the priest, that ye should be officers in the house of the LORD, for every man that is mad, and maketh himself a prophet, that thou shouldest put him in prison, and in the stocks.

Ioiadae sacerdote ut sis dux in domo Domini super omnem uirum arrepticium et prophetantem ut mittas eum in neruum

29:27 Now therefore why hast thou not reproved Jeremiah et in carcerem 29:27 et nunc quare non increpasti Hiereof Anathoth, which maketh himself a prophet to you?

29:28 For therefore he sent unto us in Babylon, saying, miam Anathothiten qui prophetat uobis 29:28 quia super

This captivity is long: build ye houses, and dwell in them; and plant gardens, and eat the fruit of them.

hoc misit ad nos in Babylonem dicens longum est aedificate domos et habitate et plantate hortos et comedite fructum eo-

29:29 And Zephaniah the priest read this letter in the ears rum 29:29 legit ergo Sophonias sacerdos librum istum in of Jeremiah the prophet.

29:30 Then came the word of the LORD unto Jeremiah, auribus Hieremiae prophetae 29:30 et factum est uerbum

saying, 29:31 Send to all them of the captivity, saying, Thus saith Domini ad Hieremiam dicens 29:31 mitte ad omnem transthe LORD concerning Shemaiah the Nehelamite; Because that Shemaiah hath prophesied unto you, and I sent him

not, and he caused you to trust in a lie: migrationem dicens haec dicit Dominus ad Semeiam Nee-

lamiten pro eo quod prophetauit uobis Semeias et ego non

29:32 Therefore thus saith the LORD; Behold, I will pun- misi eum et fecit uos confidere in mendacio 29:32 idcirco ish Shemaiah the Nehelamite, and his seed: he shall not have a man to dwell among this people; neither shall he behold the good that I will do for my people, saith the LORD; because he hath taught rebellion against the LORD.

haec dicit Dominus ecce ego uisitabo super Semeiam Neelamiten et super semen eius non erit ei uir sedens in medio populi huius et non uidebit bonum quod ego faciam populo

Prophetia Ieremiae

meo ait Dominus quia praeuaricationem locutus est aduersum Dominum

30:1 hoc uerbum quod factum est ad Hieremiam a Domino 30:1 The word that came to Jeremiah from the LORD, say-

ing,

dicens 30:2 haec dicit Dominus Deus Israhel dicens scribe 30:2 Thus speaketh the LORD God of Israel, saying, Write

thee all the words that I have spoken unto thee in a book.

tibi omnia uerba quae locutus sum ad te in libro 30:3 ecce 30:3 For, lo, the days come, saith the LORD, that I will

bring again the captivity of my people Israel and Judah, saith the LORD: and I will cause them to return to the

enim dies ueniunt dicit Dominus et conuertam conuersio- land that I gave to their fathers, and they shall possess it.

nem populi mei Israhel et Iuda ait Dominus et conuertam eos ad terram quam dedi patribus eorum et possidebunt eam 30:4 et haec uerba quae locutus est Dominus ad Israhel et 30:4 And these are the words that the LORD spake con-

cerning Israel and concerning Judah.

ad Iudam 30:5 quoniam haec dicit Dominus uocem terroris 30:5 For thus saith the LORD; We have heard a voice of

trembling, of fear, and not of peace.

audiuimus formido et non est pax 30:6 interrogate et uidete 30:6 Ask ye now, and see whether a man doth travail with

child? wherefore do I see every man with his hands on his loins, as a woman in travail, and all faces are turned into

si generat masculus quare ergo uidi omnis uiri manum super paleness? lumbum suum quasi parientis et conuersae sunt uniuersae

facies in auruginem 30:7 uae quia magna dies illa nec est 30:7 Alas! for that day is great, so that none is like it: it is even the time of Jacob’s trouble, but he shall be saved out of it.

similis eius tempusque tribulationis est Iacob et ex ipso saluabitur 30:8 et erit in die illa ait Dominus exercituum con- 30:8 For it shall come to pass in that day, saith the LORD

of hosts, that I will break his yoke from off thy neck, and will burst thy bonds, and strangers shall no more serve

teram iugum eius de collo tuo et uincula illius disrumpam et themselves of him: non dominabuntur ei amplius alieni 30:9 sed seruient Do- 30:9 But they shall serve the LORD their God, and David

their king, whom I will raise up unto them.

mino Deo suo et Dauid regi suo quem suscitabo eis 30:10 tu 30:10 Therefore fear thou not, O my servant Jacob, saith

the LORD; neither be dismayed, O Israel: for, lo, I will save thee from afar, and thy seed from the land of their captivity; and Jacob shall return, and shall be in rest, and be quiet, and none shall make him afraid.

ergo ne timeas serue meus Iacob ait Dominus neque paueas Israhel quia ecce ego saluo te de terra longinqua et semen tuum de terra captiuitatis eorum et reuertetur Iacob et quiescet et cunctis affluet et non erit quem formidet 30:11 quo- 30:11 For I am with thee, saith the LORD, to save thee:

though I make a full end of all nations whither I have scattered thee, yet I will not make a full end of thee: but I will correct thee in measure, and will not leave thee altogether unpunished.

niam tecum ego sum ait Dominus ut saluem te faciam enim consummationem in cunctis gentibus in quibus dispersi te te autem non faciam in consummationem sed castigabo te in iudicio ut non tibi uidearis innoxius 30:12 quia haec dicit 30:12 For thus saith the LORD, Thy bruise is incurable,

and thy wound is grievous.

Dominus insanabilis fractura tua pessima plaga tua 30:13 30:13 There is none to plead thy cause, that thou mayest

be bound up: thou hast no healing medicines.

non est qui iudicet iudicium tuum ad alligandum curationum utilitas non est tibi 30:14 omnes amatores tui obliti sunt tui 30:14 All thy lovers have forgotten thee; they seek thee

not; for I have wounded thee with the wound of an enemy, with the chastisement of a cruel one, for the multitude of

te non quaerent plaga enim inimici percussi te castigatione thine iniquity; because thy sins were increased. crudeli propter multitudinem iniquitatis tuae dura facta sunt peccata tua 30:15 quid clamas super contritione tua insana- 30:15 Why criest thou for thine affliction? thy sorrow is

incurable for the multitude of thine iniquity: because thy sins were increased, I have done these things unto thee.

Biblia Sacra Vulgata

bilis est dolor tuus propter multitudinem iniquitatis tuae et

30:16 Therefore all they that devour thee shall be de- dura peccata tua feci haec tibi 30:16 propterea omnes qui

voured; and all thine adversaries, every one of them, shall go into captivity; and they that spoil thee shall be a spoil,

and all that prey upon thee will I give for a prey. comedunt te deuorabuntur et uniuersi hostes tui in captiuita-

tem ducentur et qui te uastant uastabuntur cunctosque prae-

30:17 For I will restore health unto thee, and I will heal datores tuos dabo in praedam 30:17 obducam enim cicatri-

thee of thy wounds, saith the LORD; because they called thee an Outcast, saying, This is Zion, whom no man

seeketh after. cem tibi et a uulneribus tuis sanabo te dicit Dominus quia

Eiectam uocauerunt te Sion haec est quae non habebat re-

30:18 Thus saith the LORD; Behold, I will bring again the quirentem 30:18 haec dicit Dominus ecce ego conuertam captivity of Jacob’s tents, and have mercy on his dwellingplaces; and the city shall be builded upon her own heap,

and the palace shall remain after the manner thereof. conuersionem tabernaculorum Iacob et tectis eius miserebor

et aedificabitur ciuitas in excelso suo et templum iuxta ordi-

30:19 And out of them shall proceed thanksgiving and the nem suum fundabitur 30:19 et egredietur de eis laus uoxque

voice of them that make merry: and I will multiply them, and they shall not be few; I will also glorify them, and they

shall not be small. ludentium et multiplicabo eos et non inminuentur et glorifi-30:20 Their children also shall be as aforetime, and their cabo eos et non adtenuabuntur 30:20 et erunt filii eius sicut

congregation shall be established before me, and I will punish all that oppress them.

a principio et coetus eius coram me permanebit et uisitabo

30:21 And their nobles shall be of themselves, and their aduersum omnes qui tribulant eum 30:21 et erit dux eius ex

governor shall proceed from the midst of them; and I will cause him to draw near, and he shall approach unto me: for who is this that engaged his heart to approach unto me? saith the LORD.

eo et princeps de medio eius producetur et adplicabo eum et accedet ad me quis enim iste est qui adplicet cor suum ut

30:22 And ye shall be my people, and I will be your God. adpropinquet mihi ait Dominus 30:22 et eritis mihi in po-30:23 Behold, the whirlwind of the LORD goeth forth pulum et ego ero uobis in Deum 30:23 ecce turbo Domini

with fury, a continuing whirlwind: it shall fall with pain upon the head of the wicked.

furor egrediens procella ruens in capite impiorum conquie-

30:24 The fierce anger of the LORD shall not return, until scet 30:24 non auertet iram indignationis Dominus donec

he hath done it, and until he have performed the intents of his heart: in the latter days ye shall consider it.

faciat et conpleat cogitationem cordis sui in nouissimo dierum intellegetis ea

31:1 At the same time, saith the LORD, will I be the God 31:1 in tempore illo dicit Dominus ero Deus uniuersis co-

of all the families of Israel, and they shall be my people.

31:2 Thus saith the LORD, The people which were left of gnationibus Israhel et ipsi erunt mihi in populum 31:2 haec

the sword found grace in the wilderness; even Israel, when I went to cause him to rest.

dicit Dominus inuenit gratiam in deserto populus qui reman-

31:3 The LORD hath appeared of old unto me, saying, serat gladio uadet ad requiem suam Israhel 31:3 longe Do-

Yea, I have loved thee with an everlasting love: therefore with lovingkindness have I drawn thee.

minus apparuit mihi et in caritate perpetua dilexi te ideo

31:4 Again I will build thee, and thou shalt be built, O vir- adtraxi te miserans 31:4 rursumque aedificabo te et aedigin of Israel: thou shalt again be adorned with thy tabrets, and shalt go forth in the dances of them that make merry.

ficaberis uirgo Israhel adhuc ornaberis tympanis tuis et eg-

31:5 Thou shalt yet plant vines upon the mountains of redieris in choro ludentium 31:5 adhuc plantabis uineas in

Samaria: the planters shall plant, and shall eat them as common things.

montibus Samariae plantabunt plantantes et donec tempus

31:6 For there shall be a day, that the watchmen upon the ueniat non uindemiabunt 31:6 quia erit dies in qua clama-

mount Ephraim shall cry, Arise ye, and let us go up to Zion unto the LORD our God.

Prophetia Ieremiae

bunt custodes in monte Ephraim surgite et ascendamus in

Sion ad Dominum Deum nostrum 31:7 quia haec dicit Do- 31:7 For thus saith the LORD; Sing with gladness for Jacob, and shout among the chief of the nations: publish ye, praise ye, and say, O LORD, save thy people, the remnant

minus exultate in laetitia Iacob et hinnite contra caput gen- of Israel. tium personate canite et dicite salua Domine populum tuum reliquias Israhel 31:8 ecce ego adducam eos de terra aqui- 31:8 Behold, I will bring them from the north country, and

gather them from the coasts of the earth, and with them the blind and the lame, the woman with child and her that travaileth with child together: a great company shall return thither.

lonis et congregabo eos ab extremis terrae inter quos erunt caecus et claudus et praegnans et pariens simul coetus magnus reuertentium huc 31:9 in fletu uenient et in precibus 31:9 They shall come with weeping, and with supplica-

tions will I lead them: I will cause them to walk by the rivers of waters in a straight way, wherein they shall not stumble: for I am a father to Israel, and Ephraim is my firstborn.

deducam eos et adducam eos per torrentes aquarum in uia recta et non inpingent in ea quia factus sum Israheli pater et Ephraim primogenitus meus est 31:10 audite uerbum Do- 31:10 Hear the word of the LORD, O ye nations, and de-

clare it in the isles afar off, and say, He that scattered Israel will gather him, and keep him, as a shepherd doth his

mini gentes et adnuntiate insulis quae procul sunt et dicite flock. qui dispersit Israhel congregabit eum et custodiet eum sicut pastor gregem suum 31:11 redeeit enim Dominus Iacob et 31:11 For the LORD hath redeemed Jacob, and ransomed

him from the hand of him that was stronger than he.

liberauit eum de manu potentioris 31:12 et uenient et lau- 31:12 Therefore they shall come and sing in the height

of Zion, and shall flow together to the goodness of the LORD, for wheat, and for wine, and for oil, and for the young of the flock and of the herd: and their soul shall be as a watered garden; and they shall not sorrow any more at all.

dabunt in monte Sion et confluent ad bona Domini super frumento et uino et oleo et fetu pecorum et armentorum eritque anima eorum quasi hortus inriguus et ultra non esurient 31:13 tunc laetabitur uirgo in choro iuuenes et senes simul 31:13 Then shall the virgin rejoice in the dance, both

young men and old together: for I will turn their mourning into joy, and will comfort them, and make them rejoice

et conuertam luctum eorum in gaudium et consolabor eos et from their sorrow.

laetificabo a dolore suo 31:14 et inebriabo animam sacerdo- 31:14 And I will satiate the soul of the priests with fatness, and my people shall be satisfied with my goodness, saith the LORD.

tum pinguedine et populus meus bonis meis adimplebitur ait Dominus 31:15 haec dicit Dominus uox in excelso audita 31:15 Thus saith the LORD; A voice was heard in Ramah,

lamentation, and bitter weeping; Rahel weeping for her children refused to be comforted for her children, because

est lamentationis fletus et luctus Rachel plorantis filios suos they were not. et nolentis consolari super eis quia non sunt 31:16 haec di- 31:16 Thus saith the LORD; Refrain thy voice from weep-

ing, and thine eyes from tears: for thy work shall be rewarded, saith the LORD; and they shall come again from

cit Dominus quiescat uox tua a ploratu et oculi tui a lacrimis the land of the enemy. quia est merces operi tuo ait Dominus et reuertentur de terra inimici 31:17 et est spes nouissimis tuis ait Dominus et re- 31:17 And there is hope in thine end, saith the LORD, that

thy children shall come again to their own border.

uertentur filii ad terminos suos 31:18 audiens audiui Eph- 31:18 I have surely heard Ephraim bemoaning himself

thus; Thou hast chastised me, and I was chastised, as a bullock unaccustomed to the yoke: turn thou me, and I

raim transmigrantem castigasti me et eruditus sum quasi iu- shall be turned; for thou art the LORD my God. uenculus indomitus conuerte me et reuertar quia tu Dominus Deus meus 31:19 postquam enim conuertisti me egi paeni- 31:19 Surely after that I was turned, I repented; and af-

ter that I was instructed, I smote upon my thigh: I was ashamed, yea, even confounded, because I did bear the

tentiam et postquam ostendisti mihi percussi femur meum reproach of my youth.

Biblia Sacra Vulgata

confusus sum et erubui quoniam sustinui obprobrium adule-

31:20 Is Ephraim my dear son? is he a pleasant child? scentiae meae 31:20 si filius honorabilis mihi Ephraim si

for since I spake against him, I do earnestly remember him still: therefore my bowels are troubled for him; I will

surely have mercy upon him, saith the LORD. puer delicatus quia ex quo locutus sum de eo adhuc recorda-

bor eius idcirco conturbata sunt uiscera mea super eum mi-

31:21 Set thee up waymarks, make thee high heaps: set serans miserebor eius ait Dominus 31:21 statue tibi specu-

thine heart toward the highway, even the way which thou wentest: turn again, O virgin of Israel, turn again to these

thy cities. lam pone tibi amaritudines dirige cor tuum in uiam directam

in qua ambulasti reuertere uirgo Israhel reuertere ad ciuita-

31:22 How long wilt thou go about, O thou backsliding tes tuas istas 31:22 usquequo deliciis dissolueris filia uaga

daughter? for the LORD hath created a new thing in the earth, A woman shall compass a man.

quia creauit Dominus nouum super terram femina circumda-

31:23 Thus saith the LORD of hosts, the God of Israel; As bit uirum 31:23 haec dicit Dominus exercituum Deus Isra-

yet they shall use this speech in the land of Judah and in the cities thereof, when I shall bring again their captivity; The LORD bless thee, O habitation of justice, and mountain of holiness.

hel adhuc dicent uerbum istud in terra Iuda et in urbibus eius cum conuertero captiuitatem eorum benedicat tibi Dominus

31:24 And there shall dwell in Judah itself, and in all the pulchritudo iustitiae mons sanctus 31:24 et habitabunt in

cities thereof together, husbandmen, and they that go forth with flocks.

eo Iudas et omnes ciuitates eius simul agricolae et minan-

31:25 For I have satiated the weary soul, and I have re- tes greges 31:25 quia inebriaui animam lassam et omnem

plenished every sorrowful soul.

31:26 Upon this I awaked, and beheld; and my sleep was animam esurientem saturaui 31:26 ideo quasi de somno su-

sweet unto me.

31:27 Behold, the days come, saith the LORD, that I will scitatus sum et uidi et somnus meus dulcis mihi 31:27 ecce

sow the house of Israel and the house of Judah with the seed of man, and with the seed of beast.

dies ueniunt dicit Dominus et seminabo domum Israhel et

31:28 And it shall come to pass, that like as I have watched domum Iuda semine hominis et semine iumentorum 31:28 over them, to pluck up, and to break down, and to throw down, and to destroy, and to afflict; so will I watch over

them, to build, and to plant, saith the LORD. et sicut uigilaui super eos ut euellerem et demolirer et dis-

siparem et disperderem et adfligerem sic uigilabo super eos

31:29 In those days they shall say no more, The fathers ut aedificem et plantem ait Dominus 31:29 in diebus illis

have eaten a sour grape, and the children’s teeth are set on edge.

non dicent ultra patres comederunt uuam acerbam et dentes

31:30 But every one shall die for his own iniquity: every filiorum obstipuerunt 31:30 sed unusquisque in iniquitate

man that eateth the sour grape, his teeth shall be set on edge.

sua morietur omnis homo qui comederit uuam acerbam ob-

31:31 Behold, the days come, saith the LORD, that I will stupescent dentes eius 31:31 ecce dies ueniunt dicit Domi-

make a new covenant with the house of Israel, and with the house of Judah:

nus et feriam domui Israhel et domui Iuda foedus nouum

31:32 Not according to the covenant that I made with their 31:32 non secundum pactum quod pepigi cum patribus uefathers in the day that I took them by the hand to bring them out of the land of Egypt; which my covenant they brake, although I was an husband unto them, saith the LORD:

stris in die qua adprehendi manum eorum ut educerem eos de terra Aegypti pactum quod irritum fecerunt et ego domi-

31:33 But this shall be the covenant that I will make with natus sum eorum dicit Dominus 31:33 sed hoc erit pactum the house of Israel; After those days, saith the LORD, I will put my law in their inward parts, and write it in their

hearts; and will be their God, and they shall be my people. quod feriam cum domo Israhel post dies illos dicit Dominus dabo legem meam in uisceribus eorum et in corde eorum sc-

Prophetia Ieremiae

ribam eam et ero eis in Deum et ipsi erunt mihi in populum 31:34 et non docebunt ultra uir proximum suum et uir fratrem 31:34 And they shall teach no more every man his neigh-

bour, and every man his brother, saying, Know the LORD: for they shall all know me, from the least of them unto the greatest of them, saith the LORD: for I will forgive their iniquity, and I will remember their sin no more.

suum dicens cognoscite Dominum omnes enim cognoscent me a minimo eorum usque ad maximum ait Dominus quia propitiabor iniquitati eorum et peccati eorum non ero memor amplius 31:35 haec dicit Dominus qui dat solem in lumine 31:35 Thus saith the LORD, which giveth the sun for a

light by day, and the ordinances of the moon and of the stars for a light by night, which divideth the sea when the

diei ordinem lunae et stellarum in lumine noctis qui turbat waves thereof roar; The LORD of hosts is his name: mare et sonant fluctus eius Dominus exercituum nomen illi 31:36 si defecerint leges istae coram me dicit Dominus tunc 31:36 If those ordinances depart from before me, saith the

LORD, then the seed of Israel also shall cease from being a nation before me for ever.

et semen Israhel deficiet ut non sit gens coram me cunctis diebus 31:37 haec dicit Dominus si mensurari potuerint 31:37 Thus saith the LORD; If heaven above can be mea-

sured, and the foundations of the earth searched out beneath, I will also cast off all the seed of Israel for all that

caeli sursum et inuestigari fundamenta terrae deorsum et ego they have done, saith the LORD. abiciam uniuersum semen Israhel propter omnia quae fecerunt dicit Dominus 31:38 ecce dies ueniunt dicit Dominus 31:38 Behold, the days come, saith the LORD, that the

city shall be built to the LORD from the tower of Hananeel unto the gate of the corner.

et aedificabitur ciuitas Domino a turre Ananehel usque ad portam Anguli 31:39 et exibit ultra norma mensurae in con- 31:39 And the measuring line shall yet go forth over

against it upon the hill Gareb, and shall compass about to Goath.

spectu eius super collem Gareb et circuibit Goatha 31:40 et

31:40 And the whole valley of the dead bodies, and of the ashes, and all the fields unto the brook of Kidron, unto the corner of the horse gate toward the east, shall be holy unto the LORD; it shall not be plucked up, nor thrown down any more for ever.

omnem uallem cadauerum et cineris et uniuersam regionem mortis usque ad torrentem Cedron et usque ad angulum portae Equorum orientalis sanctum Domini non euelletur et non destruetur ultra in perpetuum

32:1 uerbum quod factum est ad Hieremiam a Domino in 32:1 The word that came to Jeremiah from the LORD in

the tenth year of Zedekiah king of Judah, which was the eighteenth year of Nebuchadrezzar.

anno decimo Sedeciae regis Iuda ipse est annus octauusdecimus Nabuchodonosor 32:2 tunc exercitus regis Babylo- 32:2 For then the king of Babylon’s army besieged

Jerusalem: and Jeremiah the prophet was shut up in the court of the prison, which was in the king of Judah’s

nis obsidebat Hierusalem et Hieremias propheta erat clausus house. in atrio carceris qui erat in domo regis Iuda 32:3 clauserat 32:3 For Zedekiah king of Judah had shut him up, say-

ing, Wherefore dost thou prophesy, and say, Thus saith the LORD, Behold, I will give this city into the hand of

enim eum Sedecias rex Iuda dicens quare uaticinaris dicens the king of Babylon, and he shall take it; haec dicit Dominus ecce ego dabo ciuitatem istam in manu regis Babylonis et capiet eam 32:4 et Sedecias rex Iuda non 32:4 And Zedekiah king of Judah shall not escape out of

the hand of the Chaldeans, but shall surely be delivered into the hand of the king of Babylon, and shall speak with

effugiet de manu Chaldeorum sed tradetur in manu regis Ba- him mouth to mouth, and his eyes shall behold his eyes; bylonis et loquetur os eius cum ore illius et oculi eius oculos illius uidebunt 32:5 et in Babylonem ducet Sedeciam et ibi 32:5 And he shall lead Zedekiah to Babylon, and there

shall he be until I visit him, saith the LORD: though ye fight with the Chaldeans, ye shall not prosper.

erit donec uisitem eum ait Dominus si autem dimicaueritis

Biblia Sacra Vulgata

32:6 And Jeremiah said, The word of the LORD came aduersum Chaldeos nihil prosperum habebitis 32:6 et dixit

unto me, saying, 32:7 Behold, Hanameel the son of Shallum thine uncle Hieremias factum est uerbum Domini ad me dicens 32:7 shall come unto thee saying, Buy thee my field that is in Anathoth: for the right of redemption is thine to buy it.

ecce Anamehel filius Sellum patruelis tuus ueniet ad te dicens eme tibi agrum meum qui est in Anathoth tibi enim

32:8 So Hanameel mine uncle’s son came to me in the conpetit ex propinquitate ut emas 32:8 et uenit ad me Ana-

court of the prison according to the word of the LORD, and said unto me, Buy my field, I pray thee, that is in Anathoth, which is in the country of Benjamin: for the right of inheritance is thine, and the redemption is thine; buy it for thyself. Then I knew that this was the word of the LORD.

mehel filius patrui mei secundum uerbum Domini ad uestibulum carceris et ait ad me posside agrum meum qui est in Anathoth in terra Beniamin quia tibi conpetit hereditas et tu propinquus ut possideas intellexi autem quod uerbum Do-

32:9 And I bought the field of Hanameel my uncle’s son, mini esset 32:9 et emi agrum ab Anamehel filio patrui mei

that was in Anathoth, and weighed him the money, even seventeen shekels of silver.

qui est in Anathoth et adpendi ei argentum septem stateres et

32:10 And I subscribed the evidence, and sealed it, and decem argenteos 32:10 et scripsi in libro et signaui et adhi-

took witnesses, and weighed him the money in the balances.

bui testes et adpendi argentum in statera 32:11 et accepi li-

32:11 So I took the evidence of the purchase, both that which was sealed according to the law and custom, and that which was open:

brum possessionis signatum stipulationes et rata et signa fo-

32:12 And I gave the evidence of the purchase unto Baruch rinsecus 32:12 et dedi librum possessionis Baruch filio Neri

the son of Neriah, the son of Maaseiah, in the sight of Hanameel mine uncle’s son, and in the presence of the witnesses that subscribed the book of the purchase, before all the Jews that sat in the court of the prison.

filii Maasiae in oculis Anamehel patruelis mei et in oculis testium qui scripti erant in libro emptionis in oculis omnium

32:13 And I charged Baruch before them, saying, Iudaeorum qui sedebant in atrio carceris 32:13 et praecepi 32:14 Thus saith the LORD of hosts, the God of Israel; Baruch coram eis dicens 32:14 haec dicit Dominus exerci-

Take these evidences, this evidence of the purchase, both which is sealed, and this evidence which is open; and put them in an earthen vessel, that they may continue many days.

tuum Deus Israhel sume libros istos librum emptionis hunc signatum et librum hunc qui apertus est et pones illos in uase

32:15 For thus saith the LORD of hosts, the God of Israel; fictili ut permanere possint diebus multis 32:15 haec enim

Houses and fields and vineyards shall be possessed again in this land.

dicit Dominus exercituum Deus Israhel adhuc possidebuntur

32:16 Now when I had delivered the evidence of the pur- domus et agri et uineae in terra ista 32:16 et oraui ad Domi-

chase unto Baruch the son of Neriah, I prayed unto the LORD, saying,

num postquam tradidi librum possessionis Baruch filio Neri

32:17 Ah Lord GOD! behold, thou hast made the heaven dicens 32:17 heu heu heu Domine Deus ecce tu fecisti cae-

and the earth by thy great power and stretched out arm, and there is nothing too hard for thee:

lum et terram in fortitudine tua magna et in brachio tuo ex-

32:18 Thou shewest lovingkindness unto thousands, and tento non erit tibi difficile omne uerbum 32:18 qui facis mi-

recompensest the iniquity of the fathers into the bosom of their children after them: the Great, the Mighty God, the

LORD of hosts, is his name, sericordiam in milibus et reddes iniquitatem patrum in sinu

filiorum eorum post eos fortissime magne potens Dominus

32:19 Great in counsel, and mighty in work: for thine eyes exercituum nomen tibi 32:19 magnus consilio et inconpre-

are open upon all the ways of the sons of men: to give every one according to his ways, and according to the fruit

of his doings: hensibilis cogitatu cuius oculi aperti sunt super omnes uias filiorum Adam ut reddas unicuique secundum uias suas et

Prophetia Ieremiae

secundum fructum adinuentionum eius 32:20 qui posuisti 32:20 Which hast set signs and wonders in the land of

Egypt, even unto this day, and in Israel, and among other men; and hast made thee a name, as at this day;

signa et portenta in terra Aegypti usque ad diem hanc et in Israhel et in hominibus et fecisti tibi nomen sicut est dies haec 32:21 et eduxisti populum tuum Israhel de terra Ae- 32:21 And hast brought forth thy people Israel out of the

land of Egypt with signs, and with wonders, and with a strong hand, and with a stretched out arm, and with great

gypti in signis et in portentis et in manu robusta et in bra- terror; chio extento et in terrore magno 32:22 et dedisti eis terram 32:22 And hast given them this land, which thou didst

swear to their fathers to give them, a land flowing with milk and honey;

hanc quam iurasti patribus eorum ut dares eis terram fluentem lacte et melle 32:23 et ingressi sunt et possederunt eam 32:23 And they came in, and possessed it; but they obeyed

not thy voice, neither walked in thy law; they have done nothing of all that thou commandedst them to do: there-

et non oboedierunt uoci tuae et in lege tua non ambulaue- fore thou hast caused all this evil to come upon them: runt omnia quae mandasti eis ut facerent non fecerunt et euenerunt eis omnia mala haec 32:24 ecce munitiones extruc- 32:24 Behold the mounts, they are come unto the city to

take it; and the city is given into the hand of the Chaldeans, that fight against it, because of the sword, and of the famine, and of the pestilence: and what thou hast spoken is come to pass; and, behold, thou seest it.

tae sunt aduersum ciuitatem ut capiatur et urbs data est in manu Chaldeorum qui proeliantur aduersum eam a facie gladii et famis et pestilentiae et quaecumque locutus es acciderunt ut ipse tu cernis 32:25 et tu dicis mihi Domine De- 32:25 And thou hast said unto me, O Lord GOD, Buy

thee the field for money, and take witnesses; for the city is given into the hand of the Chaldeans.

us eme agrum argento et adhibe testes cum urbs data sit in manu Chaldeorum 32:26 et factum est uerbum Domini ad 32:26 Then came the word of the LORD unto Jeremiah,

saying,

Hieremiam dicens 32:27 ecce ego Dominus Deus uniuersae 32:27 Behold, I am the LORD, the God of all flesh: is

there any thing too hard for me?

carnis numquid mihi difficile erit omne uerbum 32:28 prop- 32:28 Therefore thus saith the LORD; Behold, I will give

this city into the hand of the Chaldeans, and into the hand of Nebuchadrezzar king of Babylon, and he shall take it:

terea haec dicit Dominus ecce ego tradam ciuitatem istam in manu Chaldeorum et in manu regis Babylonis et capiet eam 32:29 et uenient Chaldei proeliantes aduersum urbem hanc et 32:29 And the Chaldeans, that fight against this city, shall

come and set fire on this city, and burn it with the houses, upon whose roofs they have offered incense unto Baal, and poured out drink offerings unto other gods, to provoke me to anger.

succendent eam igni et conburent eam et domos in quarum domatibus sacrificabant Baal et libabant diis alienis libamina ad inritandum me 32:30 erant enim filii Israhel et filii Iuda 32:30 For the children of Israel and the children of Judah

have only done evil before me from their youth: for the children of Israel have only provoked me to anger with the

iugiter facientes malum in oculis meis ab adulescentia sua work of their hands, saith the LORD. filii Israhel qui usque nunc exacerbant me in opere manuum suarum dicit Dominus 32:31 quia in furore et in indigna- 32:31 For this city hath been to me as a provocation of

mine anger and of my fury from the day that they built it even unto this day; that I should remove it from before my

tione mea facta est mihi ciuitas haec a die qua aedificaue- face, runt eam usque ad diem istam qua aufertur de conspectu meo 32:32 propter malitiam filiorum Israhel et filiorum Iuda quam 32:32 Because of all the evil of the children of Israel and of

the children of Judah, which they have done to provoke me to anger, they, their kings, their princes, their priests, and their prophets, and the men of Judah, and the inhabitants of Jerusalem.

fecerunt ad iracundiam me prouocantes ipsi et reges eorum principes eorum et sacerdotes et prophetae eorum uir Iuda

Biblia Sacra Vulgata

32:33 And they have turned unto me the back, and not the et habitatores Hierusalem 32:33 et uerterunt ad me terga et

face: though I taught them, rising up early and teaching them, yet they have not hearkened to receive instruction.

non facies cum docerem eos diluculo et erudirem et nollent

32:34 But they set their abominations in the house, which audire ut acciperent disciplinam 32:34 et posuerunt idola

is called by my name, to defile it.

sua in domo in qua inuocatum est nomen meum ut pollue-

32:35 And they built the high places of Baal, which are rent eam 32:35 et aedificauerunt excelsa Baal quae sunt in

in the valley of the son of Hinnom, to cause their sons and their daughters to pass through the fire unto Molech; which I commanded them not, neither came it into my mind, that they should do this abomination, to cause Judah to sin.

ualle filii Ennom ut initiarent filios suos et filias suas Moloch quod non mandaui eis nec ascendit in cor meum ut facerent abominationem hanc et in peccatum deducerent Iudam

32:36 And now therefore thus saith the LORD, the God of 32:36 et nunc propter ista haec dicit Dominus Deus Israhel ad

Israel, concerning this city, whereof ye say, It shall be delivered into the hand of the king of Babylon by the sword,

and by the famine, and by the pestilence; ciuitatem hanc de qua uos dicitis quod tradatur in manu re-32:37 Behold, I will gather them out of all countries, gis Babylonis in gladio et in fame et in peste 32:37 ecce ego

whither I have driven them in mine anger, and in my fury, and in great wrath; and I will bring them again unto this

place, and I will cause them to dwell safely: congregabo eos de uniuersis terris ad quas eieci eos in furore

meo et in ira mea et in indignatione grandi et reducam eos

32:38 And they shall be my people, and I will be their ad locum istum et habitare eos faciam confidenter 32:38 et

God:

32:39 And I will give them one heart, and one way, that erunt mihi in populum et ego ero eis in Deum 32:39 et dabo

they may fear me for ever, for the good of them, and of their children after them:

eis cor unum et uiam unam ut timeant me uniuersis diebus et

32:40 And I will make an everlasting covenant with them, bene sit eis et filiis eorum post eos 32:40 et feriam eis pac-

that I will not turn away from them, to do them good; but I will put my fear in their hearts, that they shall not depart

from me. tum sempiternum et non desinam eis benefacere et timorem

32:41 Yea, I will rejoice over them to do them good, and I meum dabo in corde eorum ut non recedant a me 32:41 et will plant them in this land assuredly with my whole heart and with my whole soul.

laetabor super eis cum bene eis fecero et plantabo eos in terra

32:42 For thus saith the LORD; Like as I have brought all ista in ueritate in toto corde meo et in tota anima mea 32:42

this great evil upon this people, so will I bring upon them all the good that I have promised them.

quia haec dicit Dominus sicut adduxi super populum istum omne malum hoc grande sic adducam super eos omne bo-

32:43 And fields shall be bought in this land, whereof ye num quod ego loquor ad eos 32:43 et possidebuntur agri in

say, It is desolate without man or beast; it is given into the hand of the Chaldeans.

terra ista de qua uos dicitis quod deserta sit eo quod non remanserit homo et iumentum et data sit in manu Chaldeorum

32:44 Men shall buy fields for money, and subscribe ev- 32:44 agri pecunia ementur et scribentur in libro et inprimetur

idences, and seal them, and take witnesses in the land of Benjamin, and in the places about Jerusalem, and in the cities of Judah, and in the cities of the mountains, and in the cities of the valley, and in the cities of the south: for I will cause their captivity to return, saith the LORD.

signum et testis adhibebitur in terra Beniamin et in circuitu Hierusalem in ciuitatibus Iuda et in ciuitatibus montanis et in ciuitatibus campestribus et in ciuitatibus quae ad austrum sunt quia conuertam captiuitatem eorum ait Dominus

33:1 Moreover the word of the LORD came unto Jeremiah 33:1 et factum est uerbum Domini ad Hieremiam secundo

the second time, while he was yet shut up in the court of the prison, saying,

cum adhuc clausus esset in atrio carceris dicens 33:2 haec

33:2 Thus saith the LORD the maker thereof, the LORD that formed it, to establish it; the LORD is his name;

Prophetia Ieremiae

dicit Dominus qui facturus est Dominus et formaturus illud et paraturus Dominus nomen eius 33:3 clama ad me et 33:3 Call unto me, and I will answer thee, and shew thee

great and mighty things, which thou knowest not.

exaudiam te et adnuntiabo tibi grandia et firma quae nes-

cis 33:4 quia haec dicit Dominus Deus Israhel ad domos 33:4 For thus saith the LORD, the God of Israel, concerning the houses of this city, and concerning the houses of the kings of Judah, which are thrown down by the mounts,

urbis huius et ad domos regis Iuda quae destructae sunt et and by the sword; ad munitiones et gladium 33:5 uenientium ut dimicent cum 33:5 They come to fight with the Chaldeans, but it is to fill

them with the dead bodies of men, whom I have slain in mine anger and in my fury, and for all whose wickedness

Chaldeis et impleant eas cadaueribus hominum quas per- I have hid my face from this city. cussi in furore meo et in indignatione mea abscondens faciem meam a ciuitate hac propter omnem malitiam eorum 33:6 ecce ego obducam ei cicatricem et sanitatem et curabo 33:6 Behold, I will bring it health and cure, and I will cure

them, and will reveal unto them the abundance of peace and truth.

eos et reuelabo illis deprecationem pacis et ueritatis 33:7 et

33:7 And I will cause the captivity of Judah and the captivity of Israel to return, and will build them, as at the first.

conuertam conuersionem Iuda et conuersionem Hierusalem et aedificabo eos sicut a principio 33:8 et emundabo illos 33:8 And I will cleanse them from all their iniquity,

whereby they have sinned against me; and I will pardon all their iniquities, whereby they have sinned, and whereby

ab omni iniquitate sua in qua peccauerunt mihi et propitius they have transgressed against me. ero cunctis iniquitatibus eorum in quibus deliquerunt mihi et spreuerunt me 33:9 et erit mihi in nomen et in gaudium 33:9 And it shall be to me a name of joy, a praise and an

honour before all the nations of the earth, which shall hear all the good that I do unto them: and they shall fear and tremble for all the goodness and for all the prosperity that I procure unto it.

et in laudem et in exultationem cunctis gentibus terrae quae audierint omnia bona quae ego facturus sum eis et pauebunt et turbabuntur in uniuersis bonis et in omni pace quam ego faciam ei 33:10 haec dicit Dominus adhuc audietur in loco 33:10 Thus saith the LORD; Again there shall be heard in

this place, which ye say shall be desolate without man and without beast, even in the cities of Judah, and in the streets of Jerusalem, that are desolate, without man, and without inhabitant, and without beast,

isto quem uos dicitis esse desertum eo quod non sit homo et iumentum in ciuitatibus Iuda et foris Hierusalem quae desolatae sunt absque homine et absque habitatore et abs-

que pecore 33:11 uox gaudii et uox laetitiae uox sponsi et 33:11 The voice of joy, and the voice of gladness, the voice of the bridegroom, and the voice of the bride, the voice of them that shall say, Praise the LORD of hosts: for the LORD is good; for his mercy endureth for ever: and of them that shall bring the sacrifice of praise into the house of the LORD. For I will cause to return the captivity of the land, as at the first, saith the LORD.

uox sponsae uox dicentium confitemini Domino exercituum quoniam bonus Dominus quoniam in aeternum misericordia eius et portantium uota in domum Domini reducam enim conuersionem terrae sicut a principio dicit Dominus 33:12 33:12 Thus saith the LORD of hosts; Again in this place,

which is desolate without man and without beast, and in all the cities thereof, shall be an habitation of shepherds

haec dicit Dominus exercituum adhuc erit in loco isto de- causing their flocks to lie down. serto absque homine et absque iumento et in cunctis ciuitatibus eius habitaculum pastorum accubantium gregum 33:13 33:13 In the cities of the mountains, in the cities of the

vale, and in the cities of the south, and in the land of Benjamin, and in the places about Jerusalem, and in the cities of Judah, shall the flocks pass again under the hands of him that telleth them, saith the LORD.

in ciuitatibus montuosis et in ciuitatibus campestribus et in ciuitatibus quae ad austrum sunt et in terra Beniamin et in

Biblia Sacra Vulgata

circuitu Hierusalem et in ciuitatibus Iuda adhuc transibunt

33:14 Behold, the days come, saith the LORD, that I will greges ad manum numerantis ait Dominus 33:14 ecce dies perform that good thing which I have promised unto the house of Israel and to the house of Judah.

ueniunt dicit Dominus et suscitabo uerbum bonum quod lo-

33:15 In those days, and at that time, will I cause the cutus sum ad domum Israhel et ad domum Iuda 33:15 in Branch of righteousness to grow up unto David; and he shall execute judgment and righteousness in the land.

diebus illis et in tempore illo germinare faciam Dauid ger-

33:16 In those days shall Judah be saved, and Jerusalem men iustitiae et faciet iudicium et iustitiam in terra 33:16

shall dwell safely: and this is the name wherewith she shall be called, The LORD our righteousness.

in diebus illis saluabitur Iuda et Hierusalem habitabit confidenter et hoc est quod uocabit eam Dominus iustus noster

33:17 For thus saith the LORD; David shall never want a 33:17 quia haec dicit Dominus non interibit de Dauid uir qui

man to sit upon the throne of the house of Israel;

33:18 Neither shall the priests the Levites want a man be- sedeat super thronum domus Israhel 33:18 et de sacerdoti-

fore me to offer burnt offerings, and to kindle meat offerings, and to do sacrifice continually.

bus et Leuitis non interibit uir a facie mea qui offerat holocaustomata et incendat sacrificium et caedat uictimas cunctis

33:19 And the word of the LORD came unto Jeremiah, diebus 33:19 et factum est uerbum Domini ad Hieremiam

saying,

33:20 Thus saith the LORD; If ye can break my covenant dicens 33:20 haec dicit Dominus si irritum fieri potest pac-

of the day, and my covenant of the night, and that there should not be day and night in their season;

tum meum cum die et pactum meum cum nocte ut non sit

33:21 Then may also my covenant be broken with David dies et nox in tempore suo 33:21 et pactum meum irritum

my servant, that he should not have a son to reign upon his throne; and with the Levites the priests, my ministers.

esse poterit cum Dauid seruo meo ut non sit ex eo filius qui regnet in throno eius et Leuitae et sacerdotes ministri mei

33:22 As the host of heaven cannot be numbered, neither 33:22 sicuti numerari non possunt stellae caeli et metiri ha-

the sand of the sea measured: so will I multiply the seed of David my servant, and the Levites that minister unto me.

rena maris sic multiplicabo semen Dauid serui mei et Leuitas

33:23 Moreover the word of the LORD came to Jeremiah, ministros meos 33:23 et factum est uerbum Domini ad Hie-

saying, 33:24 Considerest thou not what this people have spoken, remiam dicens 33:24 numquid non uidisti quid populus hic saying, The two families which the LORD hath chosen, he hath even cast them off? thus they have despised my

people, that they should be no more a nation before them. locutus sit dicens duae cognationes quas elegerat Dominus

abiectae sunt et populum meum despexerunt eo quod non sit

33:25 Thus saith the LORD; If my covenant be not with ultra gens coram eis 33:25 haec dicit Dominus si pactum

day and night, and if I have not appointed the ordinances of heaven and earth;

meum inter diem et noctem et leges caelo et terrae non posui

33:26 Then will I cast away the seed of Jacob and David 33:26 equidem et semen Iacob et Dauid serui mei proiciam ut

my servant, so that I will not take any of his seed to be rulers over the seed of Abraham, Isaac, and Jacob: for I will cause their captivity to return, and have mercy on them.

non adsumam de semine eius principes seminis Abraham et Isaac et Iacob reducam enim conuersionem eorum et miserebor eis

34:1 The word which came unto Jeremiah from the 34:1 uerbum quod factum est ad Hieremiam a Domino

LORD, when Nebuchadnezzar king of Babylon, and all his army, and all the kingdoms of the earth of his dominion, and all the people, fought against Jerusalem, and against all the cities thereof, saying,

quando Nabuchodonosor rex Babylonis et omnis exercitus eius uniuersaque regna terrae quae erant sub potestate ma-

Prophetia Ieremiae

nus eius et omnes populi bellabant contra Hierusalem et contra omnes urbes eius dicens 34:2 haec dicit Dominus 34:2 Thus saith the LORD, the God of Israel; Go and

speak to Zedekiah king of Judah, and tell him, Thus saith the LORD; Behold, I will give this city into the hand of

Deus Israhel uade et loquere ad Sedeciam regem Iuda et di- the king of Babylon, and he shall burn it with fire: ces ad eum haec dicit Dominus ecce ego tradam ciuitatem hanc in manu regis Babylonis et succendet eam igni 34:3 34:3 And thou shalt not escape out of his hand, but shalt

surely be taken, and delivered into his hand; and thine eyes shall behold the eyes of the king of Babylon, and he shall speak with thee mouth to mouth, and thou shalt go to Babylon.

et tu non effugies de manu eius sed conprehensione capieris et in manu eius traderis et oculi tui oculos regis Babylonis uidebunt et os eius cum ore tuo loquetur et Babylonem introibis 34:4 attamen audi uerbum Domini Sedecia rex Iuda 34:4 Yet hear the word of the LORD, O Zedekiah king of

Judah; Thus saith the LORD of thee, Thou shalt not die by the sword:

haec dicit Dominus ad te non morieris in gladio 34:5 sed

34:5 But thou shalt die in peace: and with the burnings of thy fathers, the former kings which were before thee, so shall they burn odours for thee; and they will lament thee, saying, Ah lord! for I have pronounced the word, saith the LORD.

in pace morieris et secundum conbustiones patrum tuorum regum priorum qui fuerunt ante te sic conburent te et uae domine plangent te quia uerbum ego locutus sum dicit Dominus 34:6 et locutus est Hieremias propheta ad Sedeciam 34:6 Then Jeremiah the prophet spake all these words unto

Zedekiah king of Judah in Jerusalem,

regem Iuda uniuersa uerba haec in Hierusalem 34:7 et ex- 34:7 When the king of Babylon’s army fought against

Jerusalem, and against all the cities of Judah that were left, against Lachish, and against Azekah: for these defenced

ercitus regis Babylonis pugnabat contra Hierusalem et con- cities remained of the cities of Judah. tra omnes ciuitates Iuda quae reliquae erant contra Lachis et contra Azeca haec enim supererant de ciuitatibus Iuda urbes munitae 34:8 uerbum quod factum est ad Hieremiam a Do- 34:8 This is the word that came unto Jeremiah from the

LORD, after that the king Zedekiah had made a covenant with all the people which were at Jerusalem, to proclaim

mino postquam percussit rex Sedecias foedus cum omni po- liberty unto them; pulo in Hierusalem praedicans 34:9 ut dimitteret unusquis- 34:9 That every man should let his manservant, and every

man his maidservant, being an Hebrew or an Hebrewess, go free; that none should serve himself of them, to wit, of

que seruum suum et unusquisque ancillam suam hebraeum a Jew his brother. et hebraeam liberos et nequaquam dominarentur eis id est in Iudaeo et fratre suo 34:10 audierunt ergo omnes principes 34:10 Now when all the princes, and all the people, which

had entered into the covenant, heard that every one should let his manservant, and every one his maidservant, go free, that none should serve themselves of them any more, then they obeyed, and let them go.

et uniuersus populus qui inierant pactum ut dimitteret unusquisque seruum suum et unusquisque ancillam suam liberos et ultra non dominarentur in eis audierunt igitur et dimiserunt 34:11 et conuersi sunt deinceps et retraxerunt seruos 34:11 But afterward they turned, and caused the servants

and the handmaids, whom they had let go free, to return, and brought them into subjection for servants and

et ancillas suas quos dimiserant liberos et subiugauerunt in for handmaids. famulos et in famulas 34:12 et factum est uerbum Domini 34:12 Therefore the word of the LORD came to Jeremiah

from the LORD, saying,

ad Hieremiam a Domino dicens 34:13 haec dicit Dominus 34:13 Thus saith the LORD, the God of Israel; I made a

covenant with your fathers in the day that I brought them forth out of the land of Egypt, out of the house of bond-

Deus Israhel ego percussi foedus cum patribus uestris in die men, saying, qua eduxi eos de terra Aegypti de domo seruitutis dicens

Biblia Sacra Vulgata

34:14 At the end of seven years let ye go every man his 34:14 cum conpleti fuerint septem anni dimittat unusquis-

brother an Hebrew, which hath been sold unto thee; and when he hath served thee six years, thou shalt let him go free from thee: but your fathers hearkened not unto me, neither inclined their ear.

que fratrem suum hebraeum qui uenditus est ei et seruiet tibi sex annis et dimittes eum a te liberum et non audie-

34:15 And ye were now turned, and had done right in my runt patres uestri me nec inclinauerunt aurem suam 34:15

sight, in proclaiming liberty every man to his neighbour; and ye had made a covenant before me in the house which

is called by my name: et conuersi estis uos hodie et fecistis quod rectum est in oculis meis ut praedicaretis libertatem unusquisque ad amicum suum et inistis pactum in conspectu meo in domo in qua

34:16 But ye turned and polluted my name, and caused inuocatum est nomen meum super eam 34:16 et reuersi es-

every man his servant, and every man his handmaid, whom he had set at liberty at their pleasure, to return, and brought them into subjection, to be unto you for servants and for handmaids.

tis et commaculastis nomen meum et reduxistis unusquisque seruum suum et unusquisque ancillam suam quos dimiseratis ut essent liberi et suae potestatis et subiugastis eos ut sint

34:17 Therefore thus saith the LORD; Ye have not hear- uobis serui et ancillae 34:17 propterea haec dicit Dominus

kened unto me, in proclaiming liberty, every one to his brother, and every man to his neighbour: behold, I proclaim a liberty for you, saith the LORD, to the sword, to the pestilence, and to the famine; and I will make you to be removed into all the kingdoms of the earth.

uos non audistis me ut praedicaretis libertatem unusquisque fratri suo et unusquisque amico suo ecce ego praedico libertatem ait Dominus ad gladium et pestem et famem et dabo

34:18 And I will give the men that have transgressed my uos in commotionem cunctis regnis terrae 34:18 et dabo ui-

covenant, which have not performed the words of the covenant which they had made before me, when they cut

the calf in twain, and passed between the parts thereof, ros qui praeuaricantur foedus meum et non obseruauerunt uerba foederis quibus adsensi sunt in conspectu meo uitulum quem ceciderunt in duas partes et transierunt inter di-

34:19 The princes of Judah, and the princes of Jerusalem, uisiones eius 34:19 principes Iuda et principes Hierusalem

the eunuchs, and the priests, and all the people of the land, which passed between the parts of the calf;

eunuchi et sacerdotes et omnis populus terrae qui transierunt

34:20 I will even give them into the hand of their enemies, inter diuisiones uituli 34:20 et dabo eos in manu inimico-

and into the hand of them that seek their life: and their dead bodies shall be for meat unto the fowls of the heaven,

and to the beasts of the earth. rum suorum et in manu quaerentium animam eorum et erit

morticinum eorum in escam uolucribus caeli et bestiis terrae

34:21 And Zedekiah king of Judah and his princes will I 34:21 et Sedeciam regem Iuda et principes eius dabo in manu

give into the hand of their enemies, and into the hand of them that seek their life, and into the hand of the king of

Babylon’s army, which are gone up from you. inimicorum suorum et in manu quaerentium animam eorum

et in manu exercituum regis Babylonis qui recesserunt a uo-

34:22 Behold, I will command, saith the LORD, and cause bis 34:22 ecce ego praecipio dicit Dominus et reducam eos

them to return to this city; and they shall fight against it, and take it, and burn it with fire: and I will make the cities

of Judah a desolation without an inhabitant. in ciuitatem hanc et proeliabuntur aduersum eam et capient eam et incendent igni et ciuitates Iuda dabo in solitudinem eo quod non sit habitator

35:1 The word which came unto Jeremiah from the LORD 35:1 uerbum quod factum est ad Hieremiam a Domino in in the days of Jehoiakim the son of Josiah king of Judah, saying,

diebus Ioachim filii Iosiae regis Iuda dicens 35:2 uade ad

35:2 Go unto the house of the Rechabites, and speak unto them, and bring them into the house of the LORD, into one of the chambers, and give them wine to drink.

Prophetia Ieremiae

domum Rechabitarum et loquere eis et introduces eos in domum Domini in unam exedram thesaurorum et dabis eis bibere uinum 35:3 et adsumpsi Iezoniam filium Hieremiae fi- 35:3 Then I took Jaazaniah the son of Jeremiah, the son

of Habaziniah, and his brethren, and all his sons, and the whole house of the Rechabites;

lii Absaniae et fratres eius et omnes filios eius et uniuersam domum Rechabitarum 35:4 et introduxi eos in domum Do- 35:4 And I brought them into the house of the LORD,

into the chamber of the sons of Hanan, the son of Igdaliah, a man of God, which was by the chamber of the princes, which was above the chamber of Maaseiah the son of Shallum, the keeper of the door:

mini ad gazofilacium filiorum Anan filii Hiegedeliae hominis Dei quod erat iuxta gazofilacium principum super thesaurum Maasiae filii Sellum qui erat custos uestibuli 35:5 35:5 And I set before the sons of the house of the Rech-

abites pots full of wine, and cups, and I said unto them, Drink ye wine.

et posui coram filiis domus Rechabitarum scyphos plenos uino et calices et dixi ad eos bibite uinum 35:6 qui respon- 35:6 But they said, We will drink no wine: for Jonadab the

son of Rechab our father commanded us, saying, Ye shall drink no wine, neither ye, nor your sons for ever:

derunt non bibemus uinum quia Ionadab filius Rechab pater noster praecepit nobis dicens non bibetis uinum uos et

filii uestri usque in sempiternum 35:7 et domum non aedi- 35:7 Neither shall ye build house, nor sow seed, nor plant vineyard, nor have any: but all your days ye shall dwell in tents; that ye may live many days in the land where ye be

ficabitis et sementem non seretis et uineas non plantabitis strangers. nec habebitis sed in tabernaculis habitabitis cunctis diebus uestris ut uiuatis diebus multis super faciem terrae in qua uos peregrinamini 35:8 oboediuimus ergo uoci Ionadab fi- 35:8 Thus have we obeyed the voice of Jonadab the son of

Rechab our father in all that he hath charged us, to drink no wine all our days, we, our wives, our sons, nor our

lii Rechab patris nostri in omnibus quae praecepit nobis ita daughters; ut non biberemus uinum cunctis diebus nostris nos et mulieres nostrae filii et filiae nostrae 35:9 et non aedificaremus 35:9 Nor to build houses for us to dwell in: neither have

we vineyard, nor field, nor seed:

domos ad habitandum et uineam et agrum et sementem non habuimus 35:10 sed habitauimus in tabernaculis et oboe- 35:10 But we have dwelt in tents, and have obeyed, and

done according to all that Jonadab our father commanded us.

dientes fecimus iuxta omnia quae praecepit nobis Ionadab pater noster 35:11 cum autem ascendisset Nabuchodonosor 35:11 But it came to pass, when Nebuchadrezzar king of

Babylon came up into the land, that we said, Come, and let us go to Jerusalem for fear of the army of the Chaldeans, and for fear of the army of the Syrians: so we dwell at Jerusalem.

rex Babylonis ad terram nostram diximus uenite et ingrediamur Hierusalem a facie exercitus Chaldeorum et a facie exercitus Syriae et mansimus in Hierusalem 35:12 et fac- 35:12 Then came the word of the LORD unto Jeremiah,

saying,

tum est uerbum Domini ad Hieremiam dicens 35:13 haec 35:13 Thus saith the LORD of hosts, the God of Israel; Go and tell the men of Judah and the inhabitants of Jerusalem, Will ye not receive instruction to hearken to my words?

dicit Dominus exercituum Deus Israhel uade et dic uiris Iuda saith the LORD. et habitatoribus Hierusalem numquid non recipietis disciplinam ut oboediatis uerbis meis dicit Dominus 35:14 praeua- 35:14 The words of Jonadab the son of Rechab, that he

commanded his sons not to drink wine, are performed; for unto this day they drink none, but obey their father’s commandment: notwithstanding I have spoken unto you, rising early and speaking; but ye hearkened not unto me.

luerunt sermones Ionadab filii Rechab quos praecepit filiis suis ut non biberent uinum et non biberunt usque ad diem

Biblia Sacra Vulgata

hanc quia oboedierunt praecepto patris sui ego autem locutus sum ad uos de mane consurgens et loquens et non oboedis-

35:15 I have sent also unto you all my servants the tis mihi 35:15 misique ad uos omnes seruos meos prophe-

prophets, rising up early and sending them, saying, Return ye now every man from his evil way, and amend your doings, and go not after other gods to serve them, and ye shall dwell in the land which I have given to you and to your fathers: but ye have not inclined your ear, nor hearkened unto me.

tas consurgens diluculo mittensque et dicens conuertimini unusquisque a uia sua pessima et bona facite studia uestra et nolite sequi deos alienos neque colatis eos et habitabitis in terra quam dedi uobis et patribus uestris et non inclinastis

35:16 Because the sons of Jonadab the son of Rechab aurem uestram neque audistis me 35:16 firmauerunt igitur

have performed the commandment of their father, which he commanded them; but this people hath not hearkened

unto me: filii Ionadab filii Rechab praeceptum patris sui quod praece-35:17 Therefore thus saith the LORD God of hosts, the perat eis populus autem iste non oboediuit mihi 35:17 id-

God of Israel; Behold, I will bring upon Judah and upon all the inhabitants of Jerusalem all the evil that I have pronounced against them: because I have spoken unto them, but they have not heard; and I have called unto them, but they have not answered.

circo haec dicit Dominus exercituum Deus Israhel ecce ego adduco super Iudam et super omnes habitatores Hierusalem uniuersam adflictionem quam locutus sum aduersum eos eo quod locutus sum ad illos et non audierunt uocaui illos et non

35:18 And Jeremiah said unto the house of the Rechabites, responderunt mihi 35:18 domui autem Rechabitarum dixit Thus saith the LORD of hosts, the God of Israel; Because ye have obeyed the commandment of Jonadab your father, and kept all his precepts, and done according unto all that he hath commanded you:

Hieremias haec dicit Dominus exercituum Deus Israhel pro eo quod oboedistis praecepto Ionadab patris uestri et custodistis omnia mandata eius et fecistis uniuersa quae praecepit

35:19 Therefore thus saith the LORD of hosts, the God of uobis 35:19 propterea haec dicit Dominus exercituum Deus

Israel; Jonadab the son of Rechab shall not want a man to stand before me for ever.

Israhel non deficiet uir de stirpe Ionadab filii Rechab stans in conspectu meo cunctis diebus

36:1 And it came to pass in the fourth year of Jehoiakim 36:1 et factum est in anno quarto Ioachim filii Iosiae regis

the son of Josiah king of Judah, that this word came unto Jeremiah from the LORD, saying,

Iuda factum est uerbum hoc ad Hieremiam a Domino dicens

36:2 Take thee a roll of a book, and write therein all the 36:2 tolle uolumen libri et scribes in eo omnia uerba quae lo-

words that I have spoken unto thee against Israel, and against Judah, and against all the nations, from the day I spake unto thee, from the days of Josiah, even unto this day.

cutus sum tibi aduersum Israhel et Iudam et aduersum omnes gentes a die qua locutus sum ad te ex diebus Iosiae usque ad

36:3 It may be that the house of Judah will hear all the evil diem hanc 36:3 si forte audiente domo Iuda uniuersa mala which I purpose to do unto them; that they may return every man from his evil way; that I may forgive their iniquity

and their sin. quae ego cogito facere eis reuertatur unusquisque a uia sua 36:4 Then Jeremiah called Baruch the son of Neriah: and pessima et propitius ero iniquitati et peccato eorum 36:4

Baruch wrote from the mouth of Jeremiah all the words of the LORD, which he had spoken unto him, upon a roll of

a book. uocauit ergo Hieremias Baruch filium Neriae et scripsit Ba-

ruch ex ore Hieremiae omnes sermones Domini quos locu-

36:5 And Jeremiah commanded Baruch, saying, I am shut tus est ad eum in uolumine libri 36:5 et praecepit Hiere-

up; I cannot go into the house of the LORD:

mias Baruch dicens ego clausus sum nec ualeo ingredi do-

Prophetia Ieremiae

mum Domini 36:6 ingredere ergo tu et lege de uolumine in 36:6 Therefore go thou, and read in the roll, which thou

hast written from my mouth, the words of the LORD in the ears of the people in the LORD’s house upon the fasting day: and also thou shalt read them in the ears of all Judah that come out of their cities.

quo scripsisti ex ore meo uerba Domini audiente populo in domo Domini in die ieiunii insuper et audiente uniuerso Iuda qui ueniunt de ciuitatibus suis leges eis 36:7 si forte cadat 36:7 It may be they will present their supplication before

the LORD, and will return every one from his evil way: for great is the anger and the fury that the LORD hath

oratio eorum in conspectu Domini et reuertatur unusquis- pronounced against this people. que a uia sua pessima quoniam magnus furor et indignatio quam locutus est Dominus aduersum populum hunc 36:8 36:8 And Baruch the son of Neriah did according to all

that Jeremiah the prophet commanded him, reading in the book the words of the LORD in the LORD’s house.

et fecit Baruch filius Neriae iuxta omnia quae praeceperat ei Hieremias propheta legens ex uolumine sermones Domini

in domo Domini 36:9 factum est autem in anno quinto Io- 36:9 And it came to pass in the fifth year of Jehoiakim the son of Josiah king of Judah, in the ninth month, that they proclaimed a fast before the LORD to all the people in Jerusalem, and to all the people that came from the cities of Judah unto Jerusalem.

achim filii Iosiae regis Iuda in mense nono praedicauerunt ieiunium in conspectu Domini omni populo in Hierusalem et uniuersae multitudini quae confluxerat de ciuitatibus Iuda in Hierusalem 36:10 legitque Baruch ex uolumine sermo- 36:10 Then read Baruch in the book the words of Jeremiah

in the house of the LORD, in the chamber of Gemariah the son of Shaphan the scribe, in the higher court, at the entry of the new gate of the LORD’s house, in the ears of all the people.

nes Hieremiae in domo Domini in gazofilacio Gamariae filii Saphan scribae in uestibulo superiori in introitu portae nouae domus Domini audiente omni populo 36:11 cumque audis- 36:11 When Michaiah the son of Gemariah, the son of

Shaphan, had heard out of the book all the words of the LORD,

set Micheas filius Gamariae filii Saphan omnes sermones Domini ex libro 36:12 descendit in domum regis ad gazofi- 36:12 Then he went down into the king’s house, into the

scribe’s chamber: and, lo, all the princes sat there, even Elishama the scribe, and Delaiah the son of Shemaiah, and Elnathan the son of Achbor, and Gemariah the son of Shaphan, and Zedekiah the son of Hananiah, and all the princes.

lacium scribae et ecce ibi omnes principes sedebant Elisama scriba et Dalaias filius Semeiae et Elnathan filius Achobor et Gamarias filius Saphan et Sedecias filius Ananiae et uniuersi principes 36:13 et nuntiauit eis Micheas omnia uerba 36:13 Then Michaiah declared unto them all the words

that he had heard, when Baruch read the book in the ears of the people.

quae audiuit legente Baruch ex uolumine in auribus populi 36:14 miserunt itaque omnes principes ad Baruch Iudi filium 36:14 Therefore all the princes sent Jehudi the son of

Nethaniah, the son of Shelemiah, the son of Cushi, unto Baruch, saying, Take in thine hand the roll wherein thou hast read in the ears of the people, and come. So Baruch the son of Neriah took the roll in his hand, and came unto them.

Nathaniae filii Selemiae filii Chusi dicentes uolumen ex quo legisti audiente populo sume in manu tua et ueni tulit ergo Baruch filius Neriae uolumen in manu sua et uenit ad eos 36:15 et dixerunt ad eum sede et lege haec in auribus nostris 36:15 And they said unto him, Sit down now, and read it

in our ears. So Baruch read it in their ears.

et legit Baruch in auribus eorum 36:16 igitur cum audissent 36:16 Now it came to pass, when they had heard all the

words, they were afraid both one and other, and said unto Baruch, We will surely tell the king of all these words.

omnia uerba obstipuerunt unusquisque ad proximum suum et dixerunt ad Baruch nuntiare debemus regi omnes sermones istos 36:17 et interrogauerunt eum dicentes indica nobis 36:17 And they asked Baruch, saying, Tell us now, How

didst thou write all these words at his mouth?

Biblia Sacra Vulgata

36:18 Then Baruch answered them, He pronounced all quomodo scripsisti omnes sermones istos ex ore eius 36:18

these words unto me with his mouth, and I wrote them with ink in the book.

dixit autem eis Baruch ex ore suo loquebatur quasi legens ad me omnes sermones istos et ego scribebam in uolumine

36:19 Then said the princes unto Baruch, Go, hide thee, atramento 36:19 et dixerunt principes ad Baruch uade et ab-

thou and Jeremiah; and let no man know where ye be.

36:20 And they went in to the king into the court, but they scondere tu et Hieremias et nemo sciat ubi sitis 36:20 et in-

laid up the roll in the chamber of Elishama the scribe, and told all the words in the ears of the king.

gressi sunt ad regem in atrium porro uolumen commendauerunt in gazofilacio Elisamae scribae et nuntiauerunt audiente

36:21 So the king sent Jehudi to fetch the roll: and he took rege omnes sermones 36:21 misitque rex Iudi ut sumeret it out of Elishama the scribe’s chamber. And Jehudi read it in the ears of the king, and in the ears of all the princes

which stood beside the king. uolumen qui tollens illud de gazofilacio Elisamae scribae le-

git audiente rege et uniuersis principibus qui stabant circa

36:22 Now the king sat in the winterhouse in the ninth regem 36:22 rex autem sedebat in domo hiemali in mense

month: and there was a fire on the hearth burning before him.

nono et posita erat arula coram eo plena prunis 36:23 cum-

36:23 And it came to pass, that when Jehudi had read three or four leaves, he cut it with the penknife, and cast it into the fire that was on the hearth, until all the roll was consumed in the fire that was on the hearth.

que legisset Iudi tres pagellas uel quattuor scidit illud scalpello scribae et proiecit in igne qui erat super arulam donec

36:24 Yet they were not afraid, nor rent their garments, consumeretur omne uolumen igni qui erat in arula 36:24 et

neither the king, nor any of his servants that heard all these words.

non timuerunt neque sciderunt uestimenta sua rex et omnes

36:25 Nevertheless Elnathan and Delaiah and Gemariah serui eius qui audierunt uniuersos sermones istos 36:25 ue-

had made intercession to the king that he would not burn the roll: but he would not hear them.

rumtamen Elnathan et Dalaias et Gamarias contradixerunt

36:26 But the king commanded Jerahmeel the son of Ham- regi ne conbureret librum et non audiuit eos 36:26 et prae-

melech, and Seraiah the son of Azriel, and Shelemiah the son of Abdeel, to take Baruch the scribe and Jeremiah the

prophet: but the LORD hid them. cepit rex Hieremahel filio Ammelech et Saraiae filio Ezrihel et Selemiae filio Abdehel ut conprehenderent Baruch scribam et Hieremiam prophetam abscondit autem eos Dominus

36:27 Then the word of the LORD came to Jeremiah, after 36:27 et factum est uerbum Domini ad Hieremiam postquam

that the king had burned the roll, and the words which Baruch wrote at the mouth of Jeremiah, saying,

conbuserat rex uolumen et sermones quos scripserat Baruch

36:28 Take thee again another roll, and write in it all the ex ore Hieremiae dicens 36:28 rursum tolle uolumen aliud former words that were in the first roll, which Jehoiakim the king of Judah hath burned.

et scribe in eo omnes sermones priores qui erant in uolumine

36:29 And thou shalt say to Jehoiakim king of Judah, Thus primo quod conbusit Ioachim rex Iuda 36:29 et ad Ioachim saith the LORD; Thou hast burned this roll, saying, Why hast thou written therein, saying, The king of Babylon shall certainly come and destroy this land, and shall cause to cease from thence man and beast?

regem Iuda dices haec dicit Dominus tu conbusisti uolumen illud dicens quare scripsisti in eo adnuntians festinus ueniet rex Babylonis et uastabit terram hanc et cessare faciet ex

36:30 Therefore thus saith the LORD of Jehoiakim king of illa hominem et iumentum 36:30 propterea haec dicit Do-Judah; He shall have none to sit upon the throne of David: and his dead body shall be cast out in the day to the heat,

and in the night to the frost. minus contra Ioachim regem Iuda non erit ex eo qui sedeat super solium Dauid et cadauer eius proicietur ad aestum per

Prophetia Ieremiae

diem et ad gelu per noctem 36:31 et uisitabo contra eum et 36:31 And I will punish him and his seed and his servants

for their iniquity; and I will bring upon them, and upon the inhabitants of Jerusalem, and upon the men of Judah, all the evil that I have pronounced against them; but they hearkened not.

contra semen eius et contra seruos eius iniquitates suas et adducam super eos et super habitatores Hierusalem et super uiros Iuda omne malum quod locutus sum ad eos et non audierunt 36:32 Hieremias autem tulit uolumen aliud et dedit 36:32 Then took Jeremiah another roll, and gave it to

Baruch the scribe, the son of Neriah; who wrote therein from the mouth of Jeremiah all the words of the book which Jehoiakim king of Judah had burned in the fire: and there were added besides unto them many like words.

illud Baruch filio Neriae scribae qui scripsit in eo ex ore Hieremiae omnes sermones libri quem conbuserat Ioachim rex Iuda igni et insuper additi sunt sermones multo plures quam ante fuerant

37:1 et regnauit rex Sedecias filius Iosiae pro Iechonia filio 37:1 And king Zedekiah the son of Josiah reigned instead

of Coniah the son of Jehoiakim, whom Nebuchadrezzar king of Babylon made king in the land of Judah.

Ioachim quem constituit regem Nabuchodonosor rex Babylonis in terra Iuda 37:2 et non oboediuit ipse et serui eius et 37:2 But neither he, nor his servants, nor the people of the

land, did hearken unto the words of the LORD, which he spake by the prophet Jeremiah.

populus terrae uerbis Domini quae locutus est in manu Hieremiae prophetae 37:3 et misit rex Sedecias Iuchal filium 37:3 And Zedekiah the king sent Jehucal the son of

Shelemiah and Zephaniah the son of Maaseiah the priest to the prophet Jeremiah, saying, Pray now unto the LORD

Selemiae et Sophoniam filium Maasiae sacerdotem ad Hie- our God for us. remiam prophetam dicens ora pro nobis Dominum Deum nostrum 37:4 Hieremias autem libere ambulabat in medio 37:4 Now Jeremiah came in and went out among the

people: for they had not put him into prison. Then Pharaoh’s army was come forth out of Egypt: and when the Chaldeans that besieged Jerusalem heard tidings of them, they departed from Jerusalem.

populi non enim miserant eum in custodiam carceris igitur exercitus Pharao egressus est Aegyptum et audientes Chaldei qui obsidebant Hierusalem huiuscemodi nuntium recesserunt ab Hierusalem 37:5 et factum est uerbum Domini 37:5 Then came the word of the LORD unto the prophet

Jeremiah saying,

ad Hieremiam prophetam dicens 37:6 haec dicit Dominus 37:6 Thus saith the LORD, the God of Israel; Thus shall ye say to the king of Judah, that sent you unto me to enquire of me; Behold, Pharaoh’s army, which is come forth to

Deus Israhel sic dicetis regi Iuda qui misit uos ad me ad in- help you, shall return to Egypt into their own land. terrogandum ecce exercitus Pharaonis qui egressus est uobis in auxilium reuertetur in terram suam in Aegyptum 37:7 et 37:7 And the Chaldeans shall come again, and fight

against this city, and take it, and burn it with fire.

rediint Chaldei et bellabunt contra ciuitatem hanc et capient eam et incendent igni 37:8 haec dicit Dominus nolite de- 37:8 Thus saith the LORD; Deceive not yourselves, say-

ing, The Chaldeans shall surely depart from us: for they shall not depart.

cipere animas uestras dicentes euntes abibunt et recedent a nobis Chaldei quia non abibunt 37:9 sed et si percusseri- 37:9 For though ye had smitten the whole army of the

Chaldeans that fight against you, and there remained but wounded men among them, yet should they rise up every

tis omnem exercitum Chaldeorum qui proeliantur aduersum man in his tent, and burn this city with fire. uos et derelicti fuerint ex eis aliqui uulnerati singuli de tentorio suo consurgent et incendent ciuitatem hanc igni 37:10 37:10 And it came to pass, that when the army of the

Chaldeans was broken up from Jerusalem for fear of Pharaoh’s army,

ergo cum recessisset exercitus Chaldeorum ab Hierusalem

Biblia Sacra Vulgata

37:11 Then Jeremiah went forth out of Jerusalem to go propter exercitum Pharaonis 37:11 egressus est Hieremias

into the land of Benjamin, to separate himself thence in the midst of the people.

de Hierusalem ut iret in terram Beniamin et diuideret ibi pos-

37:12 And when he was in the gate of Benjamin, a captain sessionem in conspectu ciuium 37:12 cumque peruenisset

of the ward was there, whose name was Irijah, the son of Shelemiah, the son of Hananiah; and he took Jeremiah the

prophet, saying, Thou fallest away to the Chaldeans. ad portam Beniamin erat ibi custos portae per uices nomine

Hierias filius Selemiae filii Ananiae et adprehendit Hiere-

37:13 Then said Jeremiah, It is false; I fall not away to the miam prophetam dicens ad Chaldeos profugis 37:13 et re-Chaldeans. But he hearkened not to him: so Irijah took Jeremiah, and brought him to the princes.

spondit Hieremias falsum est non fugio ad Chaldeos et non audiuit eum sed conprehendit Hierias Hieremiam et adduxit

37:14 Wherefore the princes were wroth with Jeremiah, eum ad principes 37:14 quam ob rem irati principes conand smote him, and put him in prison in the house of Jonathan the scribe: for they had made that the prison.

tra Hieremiam caesum eum miserunt in carcerem qui erat in domo Ionathan scribae ipse enim praepositus erat super car-

37:15 When Jeremiah was entered into the dungeon, and cerem 37:15 itaque ingressus est Hieremias in domum laci

into the cabins, and Jeremiah had remained there many days;

et in ergastula et sedit ibi Hieremias diebus multis 37:16

37:16 Then Zedekiah the king sent, and took him out: and the king asked him secretly in his house, and said, Is there any word from the LORD? And Jeremiah said, There is: for, said he, thou shalt be delivered into the hand of the king of Babylon.

mittens autem rex Sedecias tulit eum et interrogauit in domo sua abscondite et dixit putasne est sermo a Domino et dixit

37:17 Moreover Jeremiah said unto king Zedekiah, What Hieremias est et ait in manu regis Babylonis traderis 37:17

have I offended against thee, or against thy servants, or against this people, that ye have put me in prison?

et dixit Hieremias ad regem Sedeciam quid peccaui tibi et seruis tuis et populo tuo quia misisti me in domum carceris

37:18 Where are now your prophets which prophesied 37:18 ubi sunt prophetae uestri qui prophetabant uobis et di-

unto you, saying, The king of Babylon shall not come against you, nor against this land?

cebant non ueniet rex Babylonis super uos et super terram

37:19 Therefore hear now, I pray thee, O my lord the king: hanc 37:19 nunc ergo audi obsecro domine mi rex ualeat let my supplication, I pray thee, be accepted before thee; that thou cause me not to return to the house of Jonathan

the scribe, lest I die there. deprecatio mea in conspectu tuo et ne me remittas in do-37:20 Then Zedekiah the king commanded that they mum Ionathan scribae ne moriar ibi 37:20 praecepit ergo

should commit Jeremiah into the court of the prison, and that they should give him daily a piece of bread out of the bakers’ street, until all the bread in the city were spent. Thus Jeremiah remained in the court of the prison.

rex Sedecias ut traderetur Hieremias in uestibulo carceris et daretur ei torta panis cotidie excepto pulmento donec consumerentur omnes panes de ciuitate et mansit Hieremias in uestibulo carceris

38:1 Then Shephatiah the son of Mattan, and Gedaliah the 38:1 audiuit autem Saphatias filius Matthan et Gedelias fi-

son of Pashur, and Jucal the son of Shelemiah, and Pashur the son of Malchiah, heard the words that Jeremiah had

spoken unto all the people, saying, lius Phassur et Iuchal filius Selemiae et Phassur filius Mel-

chiae sermones quos Hieremias loquebatur ad omnem popu-

38:2 Thus saith the LORD, He that remaineth in this city lum dicens 38:2 haec dicit Dominus quicumque manserit in

shall die by the sword, by the famine, and by the pestilence: but he that goeth forth to the Chaldeans shall live;

for he shall have his life for a prey, and shall live. ciuitate hac morietur gladio et fame et peste qui autem profugerit ad Chaldeos uiuet et erit anima eius sospes et uiuens

Prophetia Ieremiae

38:3 haec dicit Dominus tradenda tradetur ciuitas haec in 38:3 Thus saith the LORD, This city shall surely be given

into the hand of the king of Babylon’s army, which shall take it.

manu exercitus regis Babylonis et capiet eam 38:4 et dixe-

38:4 Therefore the princes said unto the king, We beseech thee, let this man be put to death: for thus he weakeneth the hands of the men of war that remain in this city, and the hands of all the people, in speaking such words unto them: for this man seeketh not the welfare of this people, but the hurt.

runt principes regi rogamus ut occidatur homo iste de industria enim dissoluit manus uirorum bellantium qui remanserunt in ciuitate hac et manus uniuersi populi loquens ad eos iuxta uerba haec siquidem homo hic non quaerit pacem populi huius sed malum 38:5 et dixit rex Sedecias ecce ipse in 38:5 Then Zedekiah the king said, Behold, he is in your

hand: for the king is not he that can do any thing against you.

manibus uestris est nec enim fas est regem uobis quicquam negare 38:6 tulerunt ergo Hieremiam et proiecerunt eum in 38:6 Then took they Jeremiah, and cast him into the dun-

geon of Malchiah the son of Hammelech, that was in the court of the prison: and they let down Jeremiah with cords. And in the dungeon there was no water, but mire: so Jeremiah sunk in the mire.

lacu Melchiae filii Ammelech qui erat in uestibulo carceris et submiserunt Hieremiam in funibus et in lacum non erat aqua sed lutum descendit itaque Hieremias in caenum 38:7 38:7 Now when Ebedmelech the Ethiopian, one of the eu-

nuchs which was in the king’s house, heard that they had put Jeremiah in the dungeon; the king then sitting in the

audiuit autem Abdemelech Aethiops uir eunuchus qui erat gate of Benjamin; in domo regis quod misissent Hieremiam in lacum porro rex sedebat in porta Beniamin 38:8 et egressus est Abdemelech 38:8 Ebedmelech went forth out of the king’s house, and

spake to the king saying,

de domo regis et locutus est ad regem dicens 38:9 domine 38:9 My lord the king, these men have done evil in all that they have done to Jeremiah the prophet, whom they have cast into the dungeon; and he is like to die for hunger in the place where he is: for there is no more bread in the city.

mi rex malefecerunt uiri isti omnia quaecumque perpetrarunt contra Hieremiam prophetam mittentes eum in lacum ut moriatur ibi fame non sunt enim panes ultra in ciuitate 38:10 praecepit itaque rex Abdemelech Aethiopi dicens tolle 38:10 Then the king commanded Ebedmelech the

Ethiopian, saying, Take from hence thirty men with thee, and take up Jeremiah the prophet out of the dungeon, be-

tecum hinc triginta uiros et leua Hieremiam prophetam de fore he die. lacu antequam moriatur 38:11 adsumptis ergo Abdemelech 38:11 So Ebedmelech took the men with him, and went

into the house of the king under the treasury, and took thence old cast clouts and old rotten rags, and let them

secum uiris ingressus est domum regis quae erat sub cella- down by cords into the dungeon to Jeremiah. rio et tulit inde ueteres pannos et antiqua quae conputruerant et submisit ea ad Hieremiam in lacum per funiculos 38:12 38:12 And Ebedmelech the Ethiopian said unto Jeremiah,

Put now these old cast clouts and rotten rags under thine armholes under the cords. And Jeremiah did so.

dixitque Abdemelech Aethiops ad Hieremiam pone ueteres pannos et haec scissa et putrida sub cubitu manuum tuarum et subter funes fecit ergo Hieremias sic 38:13 et extraxerunt 38:13 So they drew up Jeremiah with cords, and took him

up out of the dungeon: and Jeremiah remained in the court of the prison.

Hieremiam funibus et eduxerunt eum de lacu mansit autem Hieremias in uestibulo carceris 38:14 et misit rex Sedecias 38:14 Then Zedekiah the king sent, and took Jeremiah the

prophet unto him into the third entry that is in the house of the LORD: and the king said unto Jeremiah, I will ask

et tulit ad se Hieremiam prophetam ad ostium tertium quod thee a thing; hide nothing from me. erat in domo Domini et dixit rex ad Hieremiam interrogo ego te sermonem ne abscondas a me aliquid 38:15 dixit au- 38:15 Then Jeremiah said unto Zedekiah, If I declare it

unto thee, wilt thou not surely put me to death? and if I give thee counsel, wilt thou not hearken unto me?

Biblia Sacra Vulgata

tem Hieremias ad Sedeciam si adnuntiauero tibi numquid non interficies me et si consilium tibi dedero non me au-

38:16 So Zedekiah the king sware secretly unto Jeremiah, dies 38:16 iurauit ergo rex Sedecias Hieremiae clam dicens

saying, As the LORD liveth, that made us this soul, I will not put thee to death, neither will I give thee into the hand

of these men that seek thy life. uiuit Dominus qui fecit nobis animam hanc si occidero te

et si tradidero te in manu uirorum istorum qui quaerunt ani-

38:17 Then said Jeremiah unto Zedekiah, Thus saith the mam tuam 38:17 et dixit Hieremias ad Sedeciam haec di-LORD, the God of hosts, the God of Israel; If thou wilt assuredly go forth unto the king of Babylon’s princes, then thy soul shall live, and this city shall not be burned with fire; and thou shalt live, and thine house:

cit Dominus exercituum Deus Israhel si profectus exieris ad principes regis Babylonis uiuet anima tua et ciuitas haec non

38:18 But if thou wilt not go forth to the king of Babylon’s succendetur igni et saluus eris tu et domus tua 38:18 si au-

princes, then shall this city be given into the hand of the Chaldeans, and they shall burn it with fire, and thou shalt

not escape out of their hand. tem non exieris ad principes regis Babylonis tradetur ciuitas

haec in manu Chaldeorum et succendent eam igni et tu non

38:19 And Zedekiah the king said unto Jeremiah, I am effugies de manu eorum 38:19 et dixit rex Sedecias ad Hie-

afraid of the Jews that are fallen to the Chaldeans, lest they deliver me into their hand, and they mock me.

remiam sollicitus sum propter Iudaeos qui transfugerunt ad Chaldeos ne forte tradar in manus eorum et inludant mihi

38:20 But Jeremiah said, They shall not deliver thee. 38:20 respondit autem Hieremias non te tradent audi quaeso

Obey, I beseech thee, the voice of the LORD, which I speak unto thee: so it shall be well unto thee, and thy soul

shall live. uocem Domini quam ego loquor ad te et bene tibi erit et uiuet 38:21 But if thou refuse to go forth, this is the word that anima tua 38:21 quod si nolueris egredi iste est sermo quem

the LORD hath shewed me:

38:22 And, behold, all the women that are left in the king ostendit mihi Dominus 38:22 ecce omnes mulieres quae re-

of Judah’s house shall be brought forth to the king of Babylon’s princes, and those women shall say, Thy friends have set thee on, and have prevailed against thee: thy feet are sunk in the mire, and they are turned away back.

manserunt in domo regis Iuda educentur ad principes regis Babylonis et ipsae dicent seduxerunt te et praeualuerunt aduersum te uiri pacifici tui demerserunt in caeno et lubrico

38:23 So they shall bring out all thy wives and thy chil- pedes tuos et recesserunt a te 38:23 et omnes uxores tuae et

dren to the Chaldeans: and thou shalt not escape out of their hand, but shalt be taken by the hand of the king of Babylon: and thou shalt cause this city to be burned with fire.

filii tui educentur ad Chaldeos et non effugies manus eorum sed in manu regis Babylonis capieris et ciuitatem hanc con-

38:24 Then said Zedekiah unto Jeremiah, Let no man buret igni 38:24 dixit ergo Sedecias ad Hieremiam nullus

know of these words, and thou shalt not die. 38:25 But if the princes hear that I have talked with thee, sciat uerba haec et non morieris 38:25 si autem audierint and they come unto thee, and say unto thee, Declare unto us now what thou hast said unto the king, hide it not from us, and we will not put thee to death; also what the king said unto thee:

principes quia locutus sum tecum et uenerint ad te et dixerint tibi indica nobis quid locutus sis cum rege ne celes nos

38:26 Then thou shalt say unto them, I presented my sup- et non te interficiemus et quid locutus est tecum rex 38:26

plication before the king, that he would not cause me to return to Jonathan’s house, to die there.

dices ad eos prostraui ego preces meas coram rege ne me re-

38:27 Then came all the princes unto Jeremiah, and asked duci iuberet in domum Ionathan et ibi morerer 38:27 uene-

him: and he told them according to all these words that the king had commanded. So they left off speaking with him;

for the matter was not perceived. runt ergo omnes principes ad Hieremiam et interrogauerunt eum et locutus est eis iuxta omnia uerba quae praeceperat

Prophetia Ieremiae

ei rex et cessauerunt ab eo nihil enim fuerat auditum 38:28 38:28 So Jeremiah abode in the court of the prison until

the day that Jerusalem was taken: and he was there when Jerusalem was taken.

mansit uero Hieremias in uestibulo carceris usque ad diem quo capta est Hierusalem et factum est ut caperetur Hierusalem

39:1 anno nono Sedeciae regis Iuda mense decimo uenit 39:1 In the ninth year of Zedekiah king of Judah, in the

tenth month, came Nebuchadrezzar king of Babylon and all his army against Jerusalem, and they besieged it.

Nabuchodonosor rex Babylonis et omnis exercitus eius ad Hierusalem et obsidebant eam 39:2 undecimo autem anno 39:2 And in the eleventh year of Zedekiah, in the fourth

month, the ninth day of the month, the city was broken up.

Sedeciae mense quarto quinta mensis aperta est ciuitas 39:3 39:3 And all the princes of the king of Babylon came in,

and sat in the middle gate, even Nergalsharezer, Samgarnebo, Sarsechim, Rabsaris, Nergalsharezer, Rabmag, with

et ingressi sunt omnes principes regis Babylonis et sederunt all the residue of the princes of the king of Babylon. in porta media Neregel Sereser Semegar Nabu Sarsachim Rabsares Neregel Sereser Rebmag et omnes reliqui principes regis Babylonis 39:4 cumque uidisset eos Sedecias rex 39:4 And it came to pass, that when Zedekiah the king of

Judah saw them, and all the men of war, then they fled, and went forth out of the city by night, by the way of the king’s garden, by the gate betwixt the two walls: and he went out the way of the plain.

Iuda et omnes uiri bellatores fugerunt et egressi sunt nocte de ciuitate per uiam horti regis et per portam quae erat inter duos muros et egressi sunt ad uiam deserti 39:5 persecu- 39:5 But the Chaldeans’ army pursued after them, and

overtook Zedekiah in the plains of Jericho: and when they had taken him, they brought him up to Nebuchadnezzar king of Babylon to Riblah in the land of Hamath, where he gave judgment upon him.

tus est autem eos exercitus Chaldeorum et conprehenderunt Sedeciam in campo solitudinis hiericuntinae et captum adduxerunt ad Nabuchodonosor regem Babylonis in Reblatha quae est in terra Emath et locutus est ad eum iudicia 39:6 39:6 Then the king of Babylon slew the sons of Zedekiah

in Riblah before his eyes: also the king of Babylon slew all the nobles of Judah.

et occidit rex Babylonis filios Sedeciae in Reblatha in oculis eius et omnes nobiles Iuda occidit rex Babylonis 39:7 ocu- 39:7 Moreover he put out Zedekiah’s eyes, and bound him

with chains, to carry him to Babylon.

los quoque Sedeciae eruit et uinxit eum conpedibus ut du-

ceretur in Babylonem 39:8 domum quoque regis et domum 39:8 And the Chaldeans burned the king’s house, and the houses of the people, with fire, and brake down the walls of Jerusalem.

uulgi succenderunt Chaldei igni et murum Hierusalem subuerterunt 39:9 et reliquias populi quae remanserunt in ciui- 39:9 Then Nebuzaradan the captain of the guard carried

away captive into Babylon the remnant of the people that remained in the city, and those that fell away, that fell to

tate et perfugas qui transfugerant ad eum et superfluos uulgi him, with the rest of the people that remained. qui remanserant transtulit Nabuzardan magister militum in

Babylonem 39:10 et de plebe pauperum qui nihil penitus 39:10 But Nebuzaradan the captain of the guard left of the poor of the people, which had nothing, in the land of Judah, and gave them vineyards and fields at the same time.

habebant dimisit Nabuzardan magister militum in terra Iuda et dedit eis uineas et cisternas in die illa 39:11 praeceperat 39:11 Now Nebuchadrezzar king of Babylon gave charge

concerning Jeremiah to Nebuzaradan the captain of the guard, saying,

autem Nabuchodonosor rex Babylonis de Hieremia Nabuzardan magistro militiae dicens 39:12 tolle illum et pone 39:12 Take him, and look well to him, and do him no

harm; but do unto him even as he shall say unto thee.

super eum oculos tuos nihilque ei mali facias sed ut uoluerit

Biblia Sacra Vulgata

39:13 So Nebuzaradan the captain of the guard sent, and sic facies ei 39:13 misit ergo Nabuzardan princeps militiae

Nebushasban, Rabsaris, and Nergalsharezer, Rabmag, and all the king of Babylon’s princes;

et Nabu et Sesban et Rabsares et Neregel et Sereser et Reb-

39:14 Even they sent, and took Jeremiah out of the court mag et omnes optimates regis Babylonis 39:14 miserunt et

of the prison, and committed him unto Gedaliah the son of Ahikam the son of Shaphan, that he should carry him

home: so he dwelt among the people. tulerunt Hieremiam de uestibulo carceris et tradiderunt eum

Godoliae filio Ahicam filii Saphan ut intraret domum et ha-

39:15 Now the word of the LORD came unto Jeremiah, bitaret in populo 39:15 ad Hieremiam autem factus fuerat

while he was shut up in the court of the prison, saying,

sermo Domini cum clausus esset in uestibulo carceris di-

39:16 Go and speak to Ebedmelech the Ethiopian, saying, cens 39:16 uade et dic Abdemelech Aethiopi dicens haec Thus saith the LORD of hosts, the God of Israel; Behold, I will bring my words upon this city for evil, and not for good; and they shall be accomplished in that day before thee.

dicit Dominus exercituum Deus Israhel ecce ego inducam sermones meos super ciuitatem hanc in malum et non in bo-

39:17 But I will deliver thee in that day, saith the LORD: num et erunt in conspectu tuo in die illa 39:17 et liberabo

and thou shalt not be given into the hand of the men of whom thou art afraid.

te in die illa ait Dominus et non traderis in manus uirorum

39:18 For I will surely deliver thee, and thou shalt not fall quos tu formidas 39:18 sed eruens liberabo te et gladio non by the sword, but thy life shall be for a prey unto thee: because thou hast put thy trust in me, saith the LORD.

cades sed erit tibi anima tua in salutem quia in me habuisti fiduciam ait Dominus

40:1 The word that came to Jeremiah from the LORD, 40:1 sermo qui factus est ad Hieremiam a Domino post-

after that Nebuzaradan the captain of the guard had let him go from Ramah, when he had taken him being bound in chains among all that were carried away captive of Jerusalem and Judah, which were carried away captive unto Babylon.

quam dimissus est a Nabuzardan magistro militiae de Rama quando tulit eum uinctum catenis in medio omnium qui migrabant de Hierusalem et Iuda et ducebantur in Babylonem

40:2 And the captain of the guard took Jeremiah, and said 40:2 tollens ergo princeps militiae Hieremiam dixit ad eum

unto him, The LORD thy God hath pronounced this evil upon this place.

Dominus Deus tuus locutus est malum hoc super locum is-

40:3 Now the LORD hath brought it, and done according tum 40:3 et adduxit et fecit Dominus sicut locutus est quia

as he hath said: because ye have sinned against the LORD, and have not obeyed his voice, therefore this thing is come

upon you. peccastis Domino et non audistis uocem eius et factus est 40:4 And now, behold, I loose thee this day from the uobis sermo hic 40:4 nunc ergo ecce solui te hodie de cate-

chains which were upon thine hand. If it seem good unto thee to come with me into Babylon, come; and I will look well unto thee: but if it seem ill unto thee to come with me into Babylon, forbear: behold, all the land is before thee: whither it seemeth good and convenient for thee to go, thither go.

nis quae sunt in manibus tuis si placet tibi ut uenias mecum in Babylonem ueni et ponam oculos meos super te si autem displicet tibi uenire mecum in Babylonem reside ecce omnis terra in conspectu tuo quod elegeris et quo placuerit

40:5 Now while he was not yet gone back, he said, Go tibi ut uadas illuc perge 40:5 et mecum noli uenire sed ha-

back also to Gedaliah the son of Ahikam the son of Shaphan, whom the king of Babylon hath made governor over the cities of Judah, and dwell with him among the people: or go wheresoever it seemeth convenient unto thee to go. So the captain of the guard gave him victuals and a reward, and let him go.

bita apud Godoliam filium Ahicam filii Saphan quem praeposuit rex Babylonis ciuitatibus Iudaeae habita ergo cum eo in medio populi uel quocumque placuerit tibi ut uadas uade dedit quoque ei magister militiae cibaria et munuscula et di-

Prophetia Ieremiae

misit eum 40:6 uenit autem Hieremias ad Godoliam filium 40:6 Then went Jeremiah unto Gedaliah the son of

Ahikam to Mizpah; and dwelt with him among the people that were left in the land.

Ahicam in Masphat et habitauit cum eo in medio populi qui relictus fuerat in terra 40:7 cum ergo audissent omnes prin- 40:7 Now when all the captains of the forces which were

in the fields, even they and their men, heard that the king of Babylon had made Gedaliah the son of Ahikam governor in the land, and had committed unto him men, and women, and children, and of the poor of the land, of them that were not carried away captive to Babylon;

cipes exercitus qui dispersi fuerant per regiones ipsi et socii eorum quod praefecisset rex Babylonis Godoliam filium Ahicam terrae et quod commendasset ei uiros et mulieres et paruulos et de pauperibus terrae qui non fuerant translati in Babylonem 40:8 uenerunt ad Godoliam in Masphat et Is- 40:8 Then they came to Gedaliah to Mizpah, even Ish-

mael the son of Nethaniah, and Johanan and Jonathan the sons of Kareah, and Seraiah the son of Tanhumeth, and the sons of Ephai the Netophathite, and Jezaniah the son of a Maachathite, they and their men.

mahel filius Nathaniae et Iohanan et Ionathan filii Caree et Sareas filius Thenoemeth et filii Offi qui erat de Nethophathi et Iezonias filius Maachathi ipsi et uiri eorum 40:9 et iurauit 40:9 And Gedaliah the son of Ahikam the son of Shaphan

sware unto them and to their men, saying, Fear not to serve the Chaldeans: dwell in the land, and serve the king of

eis Godolias filius Ahicam filii Saphan et comitibus eorum Babylon, and it shall be well with you. dicens nolite timere seruire Chaldeis habitate in terra et seruite regi Babylonis et bene erit uobis 40:10 ecce ego habito 40:10 As for me, behold, I will dwell at Mizpah, to serve

the Chaldeans, which will come unto us: but ye, gather ye wine, and summer fruits, and oil, and put them in your

in Masphat ut respondeam praecepto Chaldeorum qui mit- vessels, and dwell in your cities that ye have taken. tuntur ad nos uos autem colligite uindemiam et messem et oleum et condite in uasis uestris et manete in urbibus uestris quas tenetis 40:11 sed et omnes Iudaei qui erant in Moab 40:11 Likewise when all the Jews that were in Moab, and

among the Ammonites, and in Edom, and that were in all the countries, heard that the king of Babylon had left a remnant of Judah, and that he had set over them Gedaliah the son of Ahikam the son of Shaphan;

et in filiis Ammon et in Idumea et in uniuersis regionibus audito quod dedisset rex Babylonis reliquias in Iudaeam et quod praeposuisset super eos Godoliam filium Ahicam filii Saphan 40:12 reuersi sunt inquam omnes Iudaei de uniuer- 40:12 Even all the Jews returned out of all places whither

they were driven, and came to the land of Judah, to Gedaliah, unto Mizpah, and gathered wine and summer

sis locis ad quae profugerant et uenerunt in terram Iuda ad fruits very much. Godoliam in Masphat et collegerunt uinum et messem multam nimis 40:13 Iohanan autem filius Caree et omnes prin- 40:13 Moreover Johanan the son of Kareah, and all the

captains of the forces that were in the fields, came to Gedaliah to Mizpah,

cipes exercitus qui dispersi erant in regionibus uenerunt ad Godoliam in Masphat 40:14 et dixerunt ei scito quia Baalis 40:14 And said unto him, Dost thou certainly know that

Baalis the king of the Ammonites hath sent Ishmael the son of Nethaniah to slay thee? But Gedaliah the son of

rex filiorum Ammon misit Ismahel filium Nathaniae percu- Ahikam believed them not. tere animam tuam et non credidit eis Godolias filius Ahicam 40:15 Iohanan uero filius Caree dixit ad Godoliam seorsum in 40:15 Then Johanan the son of Kareah spake to Gedaliah

in Mizpah secretly saying, Let me go, I pray thee, and I will slay Ishmael the son of Nethaniah, and no man shall know it: wherefore should he slay thee, that all the Jews which are gathered unto thee should be scattered, and the remnant in Judah perish?

Masphat loquens ibo et percutiam Ismahel filium Nathaniae nullo sciente ne interficiat animam tuam et dissipentur omnes Iudaei qui congregati sunt ad te et peribunt reliquiae Iuda

Biblia Sacra Vulgata

40:16 But Gedaliah the son of Ahikam said unto Johanan 40:16 et ait Godolias filius Ahicam ad Iohanan filium Caree

the son of Kareah, Thou shalt not do this thing: for thou speakest falsely of Ishmael.

noli facere uerbum hoc falsum enim tu loqueris de Ismahel

41:1 Now it came to pass in the seventh month, that Ish- 41:1 et factum est in mense septimo uenit Ismahel filius

mael the son of Nethaniah the son of Elishama, of the seed royal, and the princes of the king, even ten men with him, came unto Gedaliah the son of Ahikam to Mizpah; and there they did eat bread together in Mizpah.

Nathaniae filii Elisama de semine regali et optimates regis et decem uiri cum eo ad Godoliam filium Ahicam in Mas-

41:2 Then arose Ishmael the son of Nethaniah, and the ten phat et comederunt ibi panes simul in Masphat 41:2 surmen that were with him, and smote Gedaliah the son of Ahikam the son of Shaphan with the sword, and slew him, whom the king of Babylon had made governor over the land.

rexit autem Ismahel filius Nathaniae et decem uiri qui erant cum eo et percusserunt Godoliam filium Ahicam filii Saphan gladio et interfecerunt eum quem praefecerat rex Babylonis

41:3 Ishmael also slew all the Jews that were with him, terrae 41:3 omnes quoque Iudaeos qui erant cum Godolia

even with Gedaliah, at Mizpah, and the Chaldeans that were found there, and the men of war.

in Masphat et Chaldeos qui repperti sunt ibi et uiros bella-

41:4 And it came to pass the second day after he had slain tores percussit Ismahel 41:4 secundo autem die postquam

Gedaliah, and no man knew it,

41:5 That there came certain from Shechem, from Shiloh, occiderat Godoliam nullo adhuc sciente 41:5 uenerunt uiri

and from Samaria, even fourscore men, having their beards shaven, and their clothes rent, and having cut themselves, with offerings and incense in their hand, to bring them to the house of the LORD.

de Sychem et de Silo et de Samaria octoginta uiri rasi barbam et scissis uestibus et squalentes munera et tus habebant

41:6 And Ishmael the son of Nethaniah went forth from in manu ut offerrent in domo Domini 41:6 egressus ergo

Mizpah to meet them, weeping all along as he went: and it came to pass, as he met them, he said unto them, Come

to Gedaliah the son of Ahikam. Ismahel filius Nathaniae in occursum eorum de Masphat in-

cedens et plorans ibat cum autem occurrisset eis dixit ad eos

41:7 And it was so, when they came into the midst of the uenite ad Godoliam filium Ahicam 41:7 qui cum uenissent

city, that Ishmael the son of Nethaniah slew them, and cast them into the midst of the pit, he, and the men that were

with him. ad medium ciuitatis interfecit eos Ismahel filius Nathaniae 41:8 But ten men were found among them that said unto circa medium laci ipse et uiri qui erant cum eo 41:8 decem

Ishmael, Slay us not: for we have treasures in the field, of wheat, and of barley, and of oil, and of honey. So he

forbare, and slew them not among their brethren. autem uiri repperti sunt inter eos qui dixerunt ad Ismahel noli occidere nos quia habemus thesauros in agro frumenti et hordei et olei et mellis et cessauit et non interfecit eos

41:9 Now the pit wherein Ishmael had cast all the cum fratribus suis 41:9 lacus autem in quem proiecerat Is-

dead bodies of the men, whom he had slain because of Gedaliah, was it which Asa the king had made for fear of Baasha king of Israel: and Ishmael the son of Nethaniah filled it with them that were slain.

mahel omnia cadauera uirorum quos percussit propter Godoliam ipse est quem fecit rex Asa propter Baasa regem Is-

41:10 Then Ishmael carried away captive all the residue rahel ipsum repleuit Ismahel filius Nathaniae occisis 41:10

of the people that were in Mizpah, even the king’s daughters, and all the people that remained in Mizpah, whom Nebuzaradan the captain of the guard had committed to Gedaliah the son of Ahikam: and Ishmael the son of Nethaniah carried them away captive, and departed to go over to the Ammonites.

et captiuas duxit Ismahel omnes reliquias populi qui erant in Masphat filias regis et uniuersum populum qui remanserat in Masphat quos commendarat Nabuzardan princeps militiae Godoliae filio Ahicam et cepit eos Ismahel filius Nathaniae

41:11 But when Johanan the son of Kareah, and all the et abiit ut transiret ad filios Ammon 41:11 audiuit autem

captains of the forces that were with him, heard of all the evil that Ishmael the son of Nethaniah had done,

Prophetia Ieremiae

Iohanan filius Caree et omnes principes bellatorum qui erant cum eo omne malum quod fecerat Ismahel filius Nathaniae 41:12 et adsumptis uniuersis uiris profecti sunt ut bellarent 41:12 Then they took all the men, and went to fight with

Ishmael the son of Nethaniah, and found him by the great waters that are in Gibeon.

aduersum Ismahel filium Nathaniae et inuenerunt eum ad aquas Multas quae sunt in Gabaon 41:13 cumque uidisset 41:13 Now it came to pass, that when all the people which

were with Ishmael saw Johanan the son of Kareah, and all the captains of the forces that were with him, then they

omnis populus qui erat cum Ismahel Iohanan filium Caree were glad. et uniuersos principes bellatorum qui erant cum eo laetati sunt 41:14 et reuersus est omnis populus quem ceperat Is- 41:14 So all the people that Ishmael had carried away cap-

tive from Mizpah cast about and returned, and went unto Johanan the son of Kareah.

mahel in Masphat reuersusque abiit ad Iohanan filium Caree 41:15 Ismahel autem filius Nathaniae fugit cum octo uiris a 41:15 But Ishmael the son of Nethaniah escaped from Jo-

hanan with eight men, and went to the Ammonites.

facie Iohanan et abiit ad filios Ammon 41:16 tulit ergo Ioha- 41:16 Then took Johanan the son of Kareah, and all the

captains of the forces that were with him, all the remnant of the people whom he had recovered from Ishmael the son of Nethaniah, from Mizpah, after that he had slain Gedaliah the son of Ahikam, even mighty men of war, and the women, and the children, and the eunuchs, whom he had brought again from Gibeon:

nan filius Caree et omnes principes bellatorum qui erant cum eo uniuersas reliquias uulgi quas reduxerat ab Ismahel filio Nathaniae de Masphat postquam percussit Godoliam filium Ahicam fortes uiros ad proelium et mulieres et pueros et eunuchos quos reduxerat de Gabaon 41:17 et abierunt et sede- 41:17 And they departed, and dwelt in the habitation of

Chimham, which is by Bethlehem, to go to enter into Egypt,

runt peregrinantes in Chamaam quae est iuxta Bethleem ut pergerent et introirent Aegyptum 41:18 a facie Chaldeorum 41:18 Because of the Chaldeans: for they were afraid

of them, because Ishmael the son of Nethaniah had slain Gedaliah the son of Ahikam, whom the king of Babylon

timebant enim eos quia percusserat Ismahel filius Nathaniae made governor in the land. Godoliam filium Ahicam quem praeposuerat rex Babylonis in terra Iuda

42:1 et accesserunt omnes principes bellatorum et Ioha- 42:1 Then all the captains of the forces, and Johanan the

son of Kareah, and Jezaniah the son of Hoshaiah, and all the people from the least even unto the greatest, came near,

nan filius Caree et Iezonias filius Osaiae et reliquum uulgus

a paruo usque ad magnum 42:2 dixeruntque ad Hieremiam 42:2 And said unto Jeremiah the prophet, Let, we beseech thee, our supplication be accepted before thee, and pray for us unto the LORD thy God, even for all this remnant; (for we are left but a few of many, as thine eyes do behold us:)

prophetam cadat oratio nostra in conspectu tuo et ora pro nobis ad Dominum Deum tuum pro uniuersis reliquiis istis quia derelicti sumus pauci de pluribus sicut oculi tui nos intuentur 42:3 et adnuntiet nobis Dominus Deus tuus uiam 42:3 That the LORD thy God may shew us the way

wherein we may walk, and the thing that we may do.

per quam pergamus et uerbum quod faciamus 42:4 dixit au- 42:4 Then Jeremiah the prophet said unto them, I have

heard you; behold, I will pray unto the LORD your God according to your words; and it shall come to pass, that whatsoever thing the LORD shall answer you, I will declare it unto you; I will keep nothing back from you.

tem ad eos Hieremias propheta audiui ecce ego oro ad Dominum Deum uestrum secundum uerba uestra omne uerbum quodcumque responderit mihi indicabo uobis nec celabo uos quicquam 42:5 et illi dixerunt ad Hieremiam sit Dominus 42:5 Then they said to Jeremiah, The LORD be a true and

faithful witness between us, if we do not even according to all things for the which the LORD thy God shall send thee to us.

Biblia Sacra Vulgata

inter nos testis ueritatis et fidei si non iuxta omne uerbum in

42:6 Whether it be good, or whether it be evil, we will quo miserit te Dominus Deus tuus ad nos sic faciemus 42:6

obey the voice of the LORD our God, to whom we send thee; that it may be well with us, when we obey the voice

of the LORD our God. siue bonum est siue malum uoci Domini Dei nostri ad quem

mittimus te oboediemus ut bene sit nobis cum audierimus

42:7 And it came to pass after ten days, that the word of uocem Domini Dei nostri 42:7 cum autem conpleti essent

the LORD came unto Jeremiah.

42:8 Then called he Johanan the son of Kareah, and all decem dies factum est uerbum Domini ad Hieremiam 42:8

the captains of the forces which were with him, and all the people from the least even to the greatest,

uocauitque Iohanan filium Caree et omnes principes bellatorum qui erant cum eo et uniuersum populum a minimo usque

42:9 And said unto them, Thus saith the LORD, the God of ad magnum 42:9 et dixit ad eos haec dicit Dominus Deus Israel, unto whom ye sent me to present your supplication before him;

Israhel ad quem misistis me ut prosternerem preces uestras

42:10 If ye will still abide in this land, then will I build in conspectu eius 42:10 si quiescentes manseritis in terra

you, and not pull you down, and I will plant you, and not pluck you up: for I repent me of the evil that I have done

unto you. hac aedificabo uos et non destruam plantabo et non euellam 42:11 Be not afraid of the king of Babylon, of whom ye iam enim placatus sum super malo quod feci uobis 42:11

are afraid; be not afraid of him, saith the LORD: for I am with you to save you, and to deliver you from his hand.

nolite timere a facie regis Babylonis quem uos pauidi formidatis nolite eum metuere dicit Dominus quia uobiscum sum

42:12 And I will shew mercies unto you, that he may have ego ut saluos faciam uos et eruam de manu eius 42:12 et mercy upon you, and cause you to return to your own land.

dabo uobis misericordiam et miserebor uestri et habitare uos

42:13 But if ye say, We will not dwell in this land, neither faciam in terra uestra 42:13 si autem dixeritis uos non habi-

obey the voice of the LORD your God,

tabimus in terra ista nec audiemus uocem Domini Dei nostri

42:14 Saying, No; but we will go into the land of Egypt, 42:14 dicentes nequaquam sed ad terram Aegypti pergemus

where we shall see no war, nor hear the sound of the trumpet, nor have hunger of bread; and there will we dwell:

ubi non uidebimus bellum et clangorem tubae non audiemus

42:15 And now therefore hear the word of the LORD, ye et famem non sustinebimus et ibi habitabimus 42:15 prop-

remnant of Judah; Thus saith the LORD of hosts, the God of Israel; If ye wholly set your faces to enter into Egypt,

and go to sojourn there; ter hoc nunc audite uerbum Domini reliquiae Iuda haec dicit Dominus exercituum Deus Israhel si posueritis faciem uestram ut ingrediamini Aegyptum et intraueritis ut ibi habi-

42:16 Then it shall come to pass, that the sword, which tetis 42:16 gladium quem uos formidatis ibi conprehendet

ye feared, shall overtake you there in the land of Egypt, and the famine, whereof ye were afraid, shall follow close

after you there in Egypt; and there ye shall die. uos in terra Aegypti et fames pro qua estis solliciti adherebit 42:17 So shall it be with all the men that set their faces uobis in Aegypto et ibi moriemini 42:17 omnesque uiri qui

to go into Egypt to sojourn there; they shall die by the sword, by the famine, and by the pestilence: and none of them shall remain or escape from the evil that I will bring upon them.

posuerint faciem suam ut ingrediantur Aegyptum et habitent ibi morientur gladio et fame et peste nullus de eis remanebit nec effugient a facie mali quod ego adferam super eos

42:18 For thus saith the LORD of hosts, the God of Israel; 42:18 quia haec dicit Dominus exercituum Deus Israhel sicut

As mine anger and my fury hath been poured forth upon the inhabitants of Jerusalem; so shall my fury be poured forth upon you, when ye shall enter into Egypt: and ye shall be an execration, and an astonishment, and a curse, and a reproach; and ye shall see this place no more.

conflatus est furor meus et indignatio mea super habitatores

Prophetia Ieremiae

Hierusalem sic conflabitur indignatio mea super uos cum ingressi fueritis Aegyptum et eritis in iusiurandum et in stuporem et in maledictum et in obprobrium et nequaquam ultra uidebitis locum istum 42:19 uerbum Domini super uos reli- 42:19 The LORD hath said concerning you, O ye remnant

of Judah; Go ye not into Egypt: know certainly that I have admonished you this day.

quiae Iuda nolite intrare Aegyptum scientes scietis quia obtestatus sum uobis hodie 42:20 quia decepistis animas ue- 42:20 For ye dissembled in your hearts, when ye sent me

unto the LORD your God, saying, Pray for us unto the LORD our God; and according unto all that the LORD

stras uos enim misistis me ad Dominum Deum nostrum di- our God shall say, so declare unto us, and we will do it. centes ora pro nobis ad Dominum Deum nostrum et iuxta omnia quaecumque dixerit tibi Dominus Deus noster sic adnuntia nobis et faciemus 42:21 et adnuntiaui uobis hodie 42:21 And now I have this day declared it to you; but ye

have not obeyed the voice of the LORD your God, nor any thing for the which he hath sent me unto you.

et non audistis uocem Domini Dei uestri super uniuersis pro quibus misit me ad uos 42:22 nunc ergo scientes scietis quia 42:22 Now therefore know certainly that ye shall die by

the sword, by the famine, and by the pestilence, in the place whither ye desire to go and to sojourn.

gladio et fame et peste moriemini in loco ad quem uoluistis intrare ut habitaretis ibi

43:1 factum est autem cum conplesset Hieremias loquens 43:1 And it came to pass, that when Jeremiah had made

an end of speaking unto all the people all the words of the LORD their God, for which the LORD their God had sent

ad populum uniuersos sermones Domini Dei eorum pro qui- him to them, even all these words, bus miserat eum Dominus Deus eorum ad illos omnia uerba haec 43:2 dixit Azarias filius Osaiae et Iohanan filius Ca- 43:2 Then spake Azariah the son of Hoshaiah, and Jo-

hanan the son of Kareah, and all the proud men, saying unto Jeremiah, Thou speakest falsely: the LORD our God hath not sent thee to say, Go not into Egypt to sojourn there:

ree et omnes uiri superbi dicentes ad Hieremiam mendacium tu loqueris non misit te Dominus Deus noster dicens ne ingrediamini Aegyptum ut habitetis illuc 43:3 sed Baruch fi- 43:3 But Baruch the son of Neriah setteth thee on against

us, for to deliver us into the hand of the Chaldeans, that they might put us to death, and carry us away captives into

lius Neriae incitat te aduersum nos ut tradat nos in manibus Babylon. Chaldeorum ut interficiat nos et transduci faciat in Baby-

lonem 43:4 et non audiuit Iohanan filius Caree et omnes 43:4 So Johanan the son of Kareah, and all the captains of the forces, and all the people, obeyed not the voice of the LORD, to dwell in the land of Judah.

principes bellatorum et uniuersus populus uocem Domini ut maneret in terra Iuda 43:5 sed tollens Iohanan filius Caree 43:5 But Johanan the son of Kareah, and all the captains

of the forces, took all the remnant of Judah, that were returned from all nations, whither they had been driven, to

et uniuersi principes bellatorum uniuersos reliquiarum Iuda dwell in the land of Judah; qui reuersi fuerant de cunctis gentibus ad quas fuerant ante dispersi ut habitarent in terra Iuda 43:6 uiros et mulieres et 43:6 Even men, and women, and children, and the king’s

daughters, and every person that Nebuzaradan the captain of the guard had left with Gedaliah the son of Ahikam the son of Shaphan, and Jeremiah the prophet, and Baruch the son of Neriah.

paruulos et filias regis et omnem animam quam reliquerat Nabuzardan princeps militiae cum Godolia filio Ahicam filii Saphan et Hieremiam prophetam et Baruch filium Neriae 43:7 et ingressi sunt terram Aegypti quia non oboedierunt 43:7 So they came into the land of Egypt: for they obeyed

not the voice of the LORD: thus came they even to Tahpanhes.

Biblia Sacra Vulgata

43:8 Then came the word of the LORD unto Jeremiah in uoci Domini et uenerunt usque ad Tafnas 43:8 et factus est

Tahpanhes, saying, 43:9 Take great stones in thine hand, and hide them in the sermo Domini ad Hieremiam in Tafnis dicens 43:9 sume in clay in the brickkiln, which is at the entry of Pharaoh’s house in Tahpanhes, in the sight of the men of Judah;

manu tua lapides grandes et absconde eos in crypta quae est sub muro latericio in porta domus Pharaonis in Tafnis cer-

43:10 And say unto them, Thus saith the LORD of hosts, nentibus uiris iudaeis 43:10 et dices ad eos haec dicit Do-

the God of Israel; Behold, I will send and take Nebuchadrezzar the king of Babylon, my servant, and will set his throne upon these stones that I have hid; and he shall spread his royal pavilion over them.

minus exercituum Deus Israhel ecce ego mittam et adsumam Nabuchodonosor regem Babylonis seruum meum et ponam thronum eius super lapides istos quos abscondi et statuet so-

43:11 And when he cometh, he shall smite the land of lium suum super eos 43:11 ueniensque percutiet terram Ae-

Egypt, and deliver such as are for death to death; and such as are for captivity to captivity; and such as are for the

sword to the sword. gypti quos in morte in morte et quos in captiuitate in captiui-43:12 And I will kindle a fire in the houses of the gods tate et quos in gladio in gladio 43:12 et succendet ignem in

of Egypt; and he shall burn them, and carry them away captives: and he shall array himself with the land of Egypt, as a shepherd putteth on his garment; and he shall go forth from thence in peace.

delubris deorum Aegypti et conburet ea et captiuos ducet illos et amicietur terra Aegypti sicut amicitur pastor pallio suo

43:13 He shall break also the images of Bethshemesh, that et egredietur inde in pace 43:13 et conteret statuas domus

is in the land of Egypt; and the houses of the gods of the Egyptians shall he burn with fire.

Solis quae sunt in terra Aegypti et delubra deorum Aegypti conburet igni

44:1 The word that came to Jeremiah concerning all the 44:1 uerbum quod factum est ad Hieremiam ad omnes Iu-Jews which dwell in the land of Egypt, which dwell at Migdol, and at Tahpanhes, and at Noph, and in the country

of Pathros, saying, daeos qui habitant in terra Aegypti habitantes in Magdolo et 44:2 Thus saith the LORD of hosts, the God of Israel; Ye in Tafnis et in Memphis et in terra Fatures dicens 44:2 haec

have seen all the evil that I have brought upon Jerusalem, and upon all the cities of Judah; and, behold, this day they

are a desolation, and no man dwelleth therein, dicit Dominus exercituum Deus Israhel uos uidistis omne malum istud quod adduxi super Hierusalem et super omnes urbes Iuda et ecce sunt desertae hodie et non est in eis ha-

44:3 Because of their wickedness which they have com- bitator 44:3 propter malitiam quam fecerunt ut me ad ira-

mitted to provoke me to anger, in that they went to burn incense, and to serve other gods, whom they knew not,

neither they, ye, nor your fathers. cundiam prouocarent et irent et sacrificarent et colerent deos 44:4 Howbeit I sent unto you all my servants the prophets, alienos quos nesciebant et illi et uos et patres uestri 44:4 et

rising early and sending them, saying, Oh, do not this abominable thing that I hate.

misi ad uos omnes seruos meos prophetas de nocte consurgens mittensque et dicens nolite facere uerbum abominatio-

44:5 But they hearkened not, nor inclined their ear to turn nis huius quam odi 44:5 et non audierunt nec inclinauerunt from their wickedness, to burn no incense unto other gods.

aurem suam ut conuerterentur a malis suis et non sacrifica-

44:6 Wherefore my fury and mine anger was poured forth, rent diis alienis 44:6 et conflata est indignatio mea et furor

and was kindled in the cities of Judah and in the streets of Jerusalem; and they are wasted and desolate, as at this day.

meus et succensa est in ciuitatibus Iuda et in plateis Hierusalem et uersae sunt in solitudinem et uastitatem secundum

44:7 Therefore now thus saith the LORD, the God of diem hanc 44:7 et nunc haec dicit Dominus exercituum De-

hosts, the God of Israel; Wherefore commit ye this great evil against your souls, to cut off from you man and woman, child and suckling, out of Judah, to leave you none to remain;

Prophetia Ieremiae

us Israhel quare uos facitis malum grande contra animas uestras ut intereat ex uobis uir et mulier paruulus et lactans de medio Iudae nec relinquatur uobis quicquam residuum 44:8 44:8 In that ye provoke me unto wrath with the works of

your hands, burning incense unto other gods in the land of Egypt, whither ye be gone to dwell, that ye might cut yourselves off, and that ye might be a curse and a reproach among all the nations of the earth?

prouocantes me in operibus manuum uestrarum sacrificando diis alienis in terra Aegypti in quam ingressi estis ut habitetis ibi et dispereatis et sitis in maledictionem et in obprobrium cunctis gentibus terrae 44:9 numquid obliti estis mala pat- 44:9 Have ye forgotten the wickedness of your fathers, and

the wickedness of the kings of Judah, and the wickedness of their wives, and your own wickedness, and the wickedness of your wives, which they have committed in the land of Judah, and in the streets of Jerusalem?

rum uestrorum et mala regum Iuda et mala uxorum eius et mala uestra et mala uxorum uestrarum quae fecerunt in terra Iuda et in regionibus Hierusalem 44:10 non sunt mundati 44:10 They are not humbled even unto this day, neither

have they feared, nor walked in my law, nor in my statutes, that I set before you and before your fathers.

usque ad diem hanc et non timuerunt et non ambulauerunt in lege et in praeceptis meis quae dedi coram uobis et coram patribus uestris 44:11 ideo haec dicit Dominus exerci- 44:11 Therefore thus saith the LORD of hosts, the God of

Israel; Behold, I will set my face against you for evil, and to cut off all Judah.

tuum Deus Israhel ecce ego pono faciem meam in uobis in

malum et disperdam omnem Iudam 44:12 et adsumam re- 44:12 And I will take the remnant of Judah, that have set their faces to go into the land of Egypt to sojourn there, and they shall all be consumed, and fall in the land of Egypt; they shall even be consumed by the sword and by the famine: they shall die, from the least even unto the greatest, by the sword and by the famine: and they shall be an execration, and an astonishment, and a curse, and a reproach.

liquias Iudae qui posuerunt facies suas ut ingrederentur terram Aegypti et habitarent ibi et consumentur omnes in terra Aegypti cadent in gladio et in fame consumentur a minimo usque ad maximum in gladio et in fame morientur et erunt in iusiurandum et in miraculum et in maledictionem et in obprobrium 44:13 et uisitabo habitatores terrae Aegypti sicut 44:13 For I will punish them that dwell in the land of

Egypt, as I have punished Jerusalem, by the sword, by the famine, and by the pestilence:

uisitaui super Hierusalem in gladio et in fame et in peste 44:14 et non erit qui effugiat et sit residuus de reliquiis Iu- 44:14 So that none of the remnant of Judah, which are

gone into the land of Egypt to sojourn there, shall escape or remain, that they should return into the land of Judah, to the which they have a desire to return to dwell there: for none shall return but such as shall escape.

daeorum qui uadunt ut peregrinentur in terra Aegypti et reuertantur in terram Iuda ad quam ipsi eleuant animas suas ut reuertantur et habitent ibi non reuertentur nisi qui fugerint 44:15 responderunt autem Hieremiae omnes uiri scien- 44:15 Then all the men which knew that their wives had

burned incense unto other gods, and all the women that stood by, a great multitude, even all the people that dwelt in the land of Egypt, in Pathros, answered Jeremiah, saying,

tes quod sacrificarent uxores eorum diis alienis et uniuersae mulieres quarum stabat multitudo grandis et omnis populus habitantium in terra Aegypti in Fatures dicens 44:16 sermo- 44:16 As for the word that thou hast spoken unto us in the

name of the LORD, we will not hearken unto thee.

nem quem locutus es ad nos in nomine Domini non audiemus ex te 44:17 sed facientes faciemus omne uerbum quod 44:17 But we will certainly do whatsoever thing goeth

forth out of our own mouth, to burn incense unto the queen of heaven, and to pour out drink offerings unto her, as we have done, we, and our fathers, our kings, and our princes, in the cities of Judah, and in the streets of Jerusalem: for then had we plenty of victuals, and were well, and saw no evil.

egreditur de ore nostro ut sacrificemus Reginae caeli et libe-

Biblia Sacra Vulgata

mus ei libamina sicut fecimus nos et patres nostri reges nostri et principes nostri in urbibus Iuda et in plateis Hierusalem et saturati sumus panibus et bene nobis erat malumque non

44:18 But since we left off to burn incense to the queen of uidimus 44:18 ex eo autem quo cessauimus sacrificare Re-

heaven, and to pour out drink offerings unto her, we have wanted all things, and have been consumed by the sword

and by the famine. ginae caeli et libare ei libamina indigemus omnibus et gladio 44:19 And when we burned incense to the queen of et fame consumpti sumus 44:19 quod si nos sacrificamus

heaven, and poured out drink offerings unto her, did we make her cakes to worship her, and pour out drink offer-

ings unto her, without our men? Reginae caeli et libamus ei libamina numquid sine uiris no-

stris fecimus ei placentas ad colendum eam et liba libandi

44:20 Then Jeremiah said unto all the people, to the men, 44:20 et dixit Hieremias ad omnem populum aduersum uiand to the women, and to all the people which had given him that answer, saying,

ros et aduersum mulieres et aduersum uniuersam plebem qui

44:21 The incense that ye burned in the cities of Judah, responderant ei uerbum dicens 44:21 numquid non sacrifi-

and in the streets of Jerusalem, ye, and your fathers, your kings, and your princes, and the people of the land, did not

the LORD remember them, and came it not into his mind? cium quod sacrificastis in ciuitatibus Iuda et in plateis Hierusalem uos et patres uestri reges uestri et principes uestri et populus terrae horum recordatus est Dominus et ascendit

44:22 So that the LORD could no longer bear, because super cor eius 44:22 et non poterat Dominus ultra portare

of the evil of your doings, and because of the abominations which ye have committed; therefore is your land a desolation, and an astonishment, and a curse, without an inhabitant, as at this day.

propter malitiam studiorum uestrorum et propter abominationes quas fecistis et facta est terra uestra in desolationem et in stuporem et in maledictum eo quod non sit habitator sic-

44:23 Because ye have burned incense, and because ye ut est dies haec 44:23 propterea quod sacrificaueritis idolis

have sinned against the LORD, and have not obeyed the voice of the LORD, nor walked in his law, nor in his statutes, nor in his testimonies; therefore this evil is happened unto you, as at this day.

et peccaueritis Domino et non audieritis uocem Domini et in lege et in praeceptis et in testimoniis eius non ambulaueritis idcirco euenerunt uobis mala haec sicut est dies haec

44:24 Moreover Jeremiah said unto all the people, and to 44:24 dixit autem Hieremias ad omnem populum et ad uniall the women, Hear the word of the LORD, all Judah that are in the land of Egypt:

uersas mulieres audite uerbum Domini omnis Iuda qui estis

44:25 Thus saith the LORD of hosts, the God of Israel, in terra Aegypti 44:25 haec inquit Dominus exercituum De-

saying; Ye and your wives have both spoken with your mouths, and fulfilled with your hand, saying, We will surely perform our vows that we have vowed, to burn incense to the queen of heaven, and to pour out drink offerings unto her: ye will surely accomplish your vows, and surely perform your vows.

us Israhel dicens uos et uxores uestrae locuti estis ore uestro et manibus uestris implestis dicentes faciamus uota nostra quae uouimus ut sacrificemus Reginae caeli et libemus ei li-

44:26 Therefore hear ye the word of the LORD, all Judah bamina implestis uota uestra et opere perpetrastis ea 44:26

that dwell in the land of Egypt; Behold, I have sworn by my great name, saith the LORD, that my name shall no more be named in the mouth of any man of Judah in all the land of Egypt, saying, The Lord GOD liveth.

ideo audite uerbum Domini omnis Iuda qui habitatis in terra Aegypti ecce ego iuraui in nomine meo magno ait Dominus quia nequaquam ultra nomen meum uocabitur ex ore omnis uiri iudaei dicentis uiuit Dominus Deus in omni terra Ae-

Prophetia Ieremiae

gypti 44:27 ecce ego uigilabo super eos in malum et non 44:27 Behold, I will watch over them for evil, and not for good: and all the men of Judah that are in the land of Egypt shall be consumed by the sword and by the famine, until

in bonum et consumentur omnes uiri Iuda qui sunt in terra there be an end of them. Aegypti gladio et fame donec penitus consumantur 44:28 et 44:28 Yet a small number that escape the sword shall re-

turn out of the land of Egypt into the land of Judah, and all the remnant of Judah, that are gone into the land of Egypt to sojourn there, shall know whose words shall stand, mine, or their’s.

qui fugerint gladium reuertentur de terra Aegypti in terram Iuda uiri pauci et scient omnes reliquiae Iuda ingredientium terram Aegypti ut habitent ibi cuius sermo conpleatur meus an illorum 44:29 et hoc uobis signum ait Dominus quod ui- 44:29 And this shall be a sign unto you, saith the LORD,

that I will punish you in this place, that ye may know that my words shall surely stand against you for evil:

sitem ego super uos in loco isto ut sciatis quia uere conplebuntur sermones mei contra uos in malum 44:30 haec dicit 44:30 Thus saith the LORD; Behold, I will give Pharao-

hhophra king of Egypt into the hand of his enemies, and into the hand of them that seek his life; as I gave Zedekiah king of Judah into the hand of Nebuchadrezzar king of Babylon, his enemy, and that sought his life.

Dominus ecce ego tradam Pharaonem Efree regem Aegypti in manu inimicorum eius et in manu quaerentium animam illius sicut tradidi Sedeciam regem Iuda in manu Nabuchodonosor regis Babylonis inimici sui et quaerentis animam eius

45:1 uerbum quod locutus est Hieremias propheta ad Ba- 45:1 The word that Jeremiah the prophet spake unto

Baruch the son of Neriah, when he had written these words in a book at the mouth of Jeremiah, in the fourth year of

ruch filium Neri cum scripsisset uerba haec in libro de ore Jehoiakim the son of Josiah king of Judah, saying, Hieremiae anno quarto Ioachim filii Iosiae regis Iuda dicens 45:2 haec dicit Dominus Deus Israhel ad te Baruch 45:3 dix- 45:2 Thus saith the LORD, the God of Israel, unto thee, O

Baruch:

45:3 Thou didst say, Woe is me now! for the LORD hath added grief to my sorrow; I fainted in my sighing, and I find no rest.

isti uae misero mihi quoniam addidit Dominus dolorem dolori meo laboraui in gemitu meo et requiem non inueni 45:4 45:4 Thus shalt thou say unto him, The LORD saith thus;

Behold, that which I have built will I break down, and that which I have planted I will pluck up, even this whole land.

haec dices ad eum sic dicit Dominus ecce quos aedificaui ego destruo et quos plantaui ego euello et uniuersam terram hanc 45:5 et tu quaeris tibi grandia noli quaerere quia ecce 45:5 And seekest thou great things for thyself? seek them

not: for, behold, I will bring evil upon all flesh, saith the LORD: but thy life will I give unto thee for a prey in all

ego adducam malum super omnem carnem ait Dominus et places whither thou goest. dabo tibi animam tuam in salutem in omnibus locis ad quaecumque perrexeris

46:1 quod factum est uerbum Domini ad Hieremiam pro- 46:1 The word of the LORD which came to Jeremiah the

prophet against the Gentiles;

phetam contra gentes 46:2 ad Aegyptum aduersum exer- 46:2 Against Egypt, against the army of Pharaohnecho

king of Egypt, which was by the river Euphrates in Carchemish, which Nebuchadrezzar king of Babylon smote in the fourth year of Jehoiakim the son of Josiah king of Judah.

citum Pharaonis Nechao regis Aegypti qui erat iuxta flumen Eufraten in Charchamis quem percussit Nabuchodonosor rex Babylonis in quarto anno Ioachim filii Iosiae regis Iuda 46:3 praeparate scutum et clypeum et procedite ad bel- 46:3 Order ye the buckler and shield, and draw near to

battle.

lum 46:4 iungite equos et ascendite equites state in galeis 46:4 Harness the horses; and get up, ye horsemen, and

stand forth with your helmets; furbish the spears, and put on the brigandines.

Biblia Sacra Vulgata

46:5 Wherefore have I seen them dismayed and turned polite lanceas induite uos loricis 46:5 quid igitur uidi ipsos

away back? and their mighty ones are beaten down, and are fled apace, and look not back: for fear was round

about, saith the LORD. pauidos et terga uertentes fortes eorum caesos fugerunt con-46:6 Let not the swift flee away, nor the mighty man es- citi nec respexerunt terror undique ait Dominus 46:6 non

cape; they shall stumble, and fall toward the north by the river Euphrates.

fugiat uelox nec saluari se putet fortis ad aquilonem iuxta

46:7 Who is this that cometh up as a flood, whose waters flumen Eufraten uicti sunt et ruerunt 46:7 quis est iste qui

are moved as the rivers?

quasi flumen ascendit et ueluti fluuiorum intumescunt gur-

46:8 Egypt riseth up like a flood, and his waters are moved gites eius 46:8 Aegyptus fluminis instar ascendet et uelut

like the rivers; and he saith, I will go up, and will cover the earth; I will destroy the city and the inhabitants thereof.

flumina mouebuntur fluctus eius et dicet ascendens operiam

46:9 Come up, ye horses; and rage, ye chariots; and let the terram perdam ciuitatem et habitatores eius 46:9 ascendite

mighty men come forth; the Ethiopians and the Libyans, that handle the shield; and the Lydians, that handle and

bend the bow. equos et exultate in curribus et procedant fortes Aethiopia et

Lybies tenentes scutum et Lydii arripientes et iacientes sa-

46:10 For this is the day of the Lord GOD of hosts, a day gittas 46:10 dies autem ille Domini Dei exercituum dies ul-

of vengeance, that he may avenge him of his adversaries: and the sword shall devour, and it shall be satiate and made drunk with their blood: for the Lord GOD of hosts hath a sacrifice in the north country by the river Euphrates.

tionis ut sumat uindictam de inimicis suis deuorabit gladius et saturabitur et inebriabitur sanguine eorum uictima enim Domini exercituum in terra aquilonis iuxta flumen Eufraten

46:11 Go up into Gilead, and take balm, O virgin, the 46:11 ascende in Galaad et tolle resinam uirgo filia Aegypti

daughter of Egypt: in vain shalt thou use many medicines; for thou shalt not be cured.

frustra multiplicas medicamina sanitas non erit tibi 46:12

46:12 The nations have heard of thy shame, and thy cry hath filled the land: for the mighty man hath stumbled against the mighty, and they are fallen both together.

audierunt gentes ignominiam tuam et ululatus tuus repleuit terram quia fortis inpegit in fortem ambo pariter conciderunt

46:13 The word that the LORD spake to Jeremiah the 46:13 uerbum quod locutus est Dominus ad Hieremiam pro-

prophet, how Nebuchadrezzar king of Babylon should come and smite the land of Egypt.

phetam super eo quod uenturus esset Nabuchodonosor rex

46:14 Declare ye in Egypt, and publish in Migdol, and Babylonis et percussurus terram Aegypti 46:14 adnuntiate

publish in Noph and in Tahpanhes: say ye, Stand fast, and prepare thee; for the sword shall devour round about thee.

Aegypto et auditum facite Magdolo et resonet in Memphis et in Tafnis dicite sta et praepara te quia deuorauit gladius ea

46:15 Why are thy valiant men swept away? they stood quae per circuitum tuum sunt 46:15 quare conputruit for-

not, because the LORD did drive them.

46:16 He made many to fall, yea, one fell upon another: tis tuus non stetit quoniam Dominus subuertit eum 46:16

and they said, Arise, and let us go again to our own people, and to the land of our nativity, from the oppressing sword.

multiplicauit ruentes ceciditque uir ad proximum suum et dicent surge et reuertamur ad populum nostrum et ad ter-

46:17 They did cry there, Pharaoh king of Egypt is but a ram natiuitatis nostrae a facie gladii columbae 46:17 uo-

noise; he hath passed the time appointed.

cate nomen Pharao regis Aegypti Tumultum adduxit tem-

46:18 As I live, saith the King, whose name is the LORD pus 46:18 uiuo ego inquit Rex Dominus exercituum nomen

of hosts, Surely as Tabor is among the mountains, and as Carmel by the sea, so shall he come.

eius quoniam sicut Thabor in montibus et sicut Carmelus

46:19 O thou daughter dwelling in Egypt, furnish thyself in mari ueniet 46:19 uasa transmigrationis fac tibi habitat-

to go into captivity: for Noph shall be waste and desolate without an inhabitant.

Prophetia Ieremiae

rix filia Aegypti quia Memphis in solitudinem erit et deseretur inhabitabilis 46:20 uitula eligans atque formonsa Ae- 46:20 Egypt is like a very fair heifer, but destruction

cometh; it cometh out of the north.

gyptus stimulator ab aquilone ueniet ei 46:21 mercennarii 46:21 Also her hired men are in the midst of her like fat-

ted bullocks; for they also are turned back, and are fled away together: they did not stand, because the day of their calamity was come upon them, and the time of their visitation.

quoque eius qui uersabantur in medio eius quasi uituli saginati uersi sunt et fugerunt simul nec stare potuerunt quia dies interfectionis eorum uenit super eos tempus uisitationis eorum 46:22 uox eius quasi aeris sonabit quoniam cum ex- 46:22 The voice thereof shall go like a serpent; for they

shall march with an army, and come against her with axes, as hewers of wood.

ercitu properabunt et cum securibus uenient ei quasi ligna caedentes 46:23 succiderunt saltum eius ait Dominus qui 46:23 They shall cut down her forest, saith the LORD,

though it cannot be searched; because they are more than the grasshoppers, and are innumerable.

supputari non potest multiplicati sunt super lucustas et non est eis numerus 46:24 confusa est filia Aegypti et tradita 46:24 The daughter of Egypt shall be confounded; she

shall be delivered into the hand of the people of the north.

in manu populi aquilonis 46:25 dixit Dominus exercituum 46:25 The LORD of hosts, the God of Israel, saith; Be-

hold, I will punish the multitude of No, and Pharaoh, and Egypt, with their gods, and their kings; even Pharaoh, and

Deus Israhel ecce ego uisitabo super tumultum Alexandriae all them that trust in him: et super Pharao et super Aegyptum et super deos eius et super reges eius et super Pharao et super eos qui confidunt in

eo 46:26 et dabo eos in manu quaerentium animam eorum 46:26 And I will deliver them into the hand of those that seek their lives, and into the hand of Nebuchadrezzar king of Babylon, and into the hand of his servants: and afterward it shall be inhabited, as in the days of old, saith the LORD.

et in manu Nabuchodonosor regis Babylonis et in manu seruorum eius et post haec habitabitur sicut diebus pristinis ait Dominus 46:27 et tu ne timeas serue meus Iacob et ne pa- 46:27 But fear not thou, O my servant Jacob, and be not

dismayed, O Israel: for, behold, I will save thee from afar off, and thy seed from the land of their captivity; and Jacob shall return, and be in rest and at ease, and none shall make him afraid.

ueas Israhel quia ecce ego saluum te faciam de longinquo et semen tuum de terra captiuitatis suae et reuertetur Iacob et quiescet et prosperabitur et non erit qui exterreat eum 46:28 46:28 Fear thou not, O Jacob my servant, saith the LORD:

for I am with thee; for I will make a full end of all the nations whither I have driven thee: but I will not make a full end of thee, but correct thee in measure; yet will I not leave thee wholly unpunished.

et tu noli timere serue meus Iacob ait Dominus quia tecum ego sum quia consumam ego cunctas gentes ad quas eieci te te uero non consumam sed castigabo te in iudicio nec quasi innocenti parcam tibi

47:1 quod factum est uerbum Domini ad Hieremiam pro- 47:1 The word of the LORD that came to Jeremiah the

prophet against the Philistines, before that Pharaoh smote Gaza.

phetam contra Palestinos antequam percuteret Pharao Gazam 47:2 haec dicit Dominus ecce aquae ascendunt ab 47:2 Thus saith the LORD; Behold, waters rise up out

of the north, and shall be an overflowing flood, and shall overflow the land, and all that is therein; the city, and them that dwell therein: then the men shall cry, and all the inhabitants of the land shall howl.

aquilone et erunt quasi torrens inundans et operient terram et plenitudinem eius urbem et habitatores eius clamabunt homines et ululabit omnis habitator terrae 47:3 ab strepitu 47:3 At the noise of the stamping of the hoofs of his strong

horses, at the rushing of his chariots, and at the rumbling of his wheels, the fathers shall not look back to their chil-

pompae armorum et bellatorum eius a commotione quadri- dren for feebleness of hands;

Biblia Sacra Vulgata

garum eius et multitudine rotarum illius non respexerunt pa-

47:4 Because of the day that cometh to spoil all the tres filios manibus dissolutis 47:4 pro aduentu diei in quo

Philistines, and to cut off from Tyrus and Zidon every helper that remaineth: for the LORD will spoil the

Philistines, the remnant of the country of Caphtor. uastabuntur omnes Philisthim et dissipabitur Tyrus et Sidon

cum omnibus reliquis auxiliis suis depopulatus est enim Do-

47:5 Baldness is come upon Gaza; Ashkelon is cut off with minus Palestinos reliquias insulae Cappadociae 47:5 uenit

the remnant of their valley: how long wilt thou cut thyself?

caluitium super Gazam conticuit Ascalon et reliquiae ual-

47:6 O thou sword of the LORD, how long will it be ere lis earum usquequo concideris 47:6 o mucro Domini us-

thou be quiet? put up thyself into thy scabbard, rest, and be still.

quequo non quiescis ingredere in uaginam tuam refrigerare

47:7 How can it be quiet, seeing the LORD hath given it a et sile 47:7 quomodo quiescet cum Dominus praeceperit ei

charge against Ashkelon, and against the sea shore? there hath he appointed it.

aduersus Ascalonem et aduersus maritimas eius regiones ibique condixerit illi

48:1 Against Moab thus saith the LORD of hosts, the God 48:1 ad Moab haec dicit Dominus exercituum Deus Israof Israel; Woe unto Nebo! for it is spoiled: Kiriathaim is confounded and taken: Misgab is confounded and dis-

mayed. hel uae super Nabo quoniam uastata est et confusa capta est 48:2 There shall be no more praise of Moab: in Heshbon Cariathaim confusa est fortis et tremuit 48:2 non est ultra

they have devised evil against it; come, and let us cut it off from being a nation. Also thou shalt be cut down, O

Madmen; the sword shall pursue thee. exultatio in Moab contra Esebon cogitauerunt malum uenite

et disperdamus eam de gente ergo silens conticesces seque-

48:3 A voice of crying shall be from Horonaim, spoiling turque te gladius 48:3 uox clamoris de Oronaim uastitas et

and great destruction.

48:4 Moab is destroyed; her little ones have caused a cry contritio magna 48:4 contrita est Moab adnuntiate clamo-

to be heard. 48:5 For in the going up of Luhith continual weeping shall rem paruulis eius 48:5 per ascensum enim Luaith plorans go up; for in the going down of Horonaim the enemies have heard a cry of destruction.

ascendet in fletu quoniam in descensu Oronaim hostes ulu-

48:6 Flee, save your lives, and be like the heath in the latum contritionis audierunt 48:6 fugite saluate animas ue-

wilderness.

48:7 For because thou hast trusted in thy works and in stras et eritis quasi myrice in deserto 48:7 pro eo enim quod

thy treasures, thou shalt also be taken: and Chemosh shall go forth into captivity with his priests and his princes to-

gether. habuisti fiduciam in munitionibus tuis et in thesauris tuis tu

quoque capieris et ibit Chamos in transmigrationem sacerdo-

48:8 And the spoiler shall come upon every city, and no tes eius et principes eius simul 48:8 et ueniet praedo ad om-

city shall escape: the valley also shall perish, and the plain shall be destroyed, as the LORD hath spoken.

nem urbem et urbs nulla saluabitur et peribit uallis et dissi-

48:9 Give wings unto Moab, that it may flee and get away: pabuntur campestria quoniam dixit Dominus 48:9 date flo-

for the cities thereof shall be desolate, without any to dwell therein.

rem Moab quia floriens egredietur et ciuitates eius desertae

48:10 Cursed be he that doeth the work of the LORD de- erunt et inhabitabiles 48:10 maledictus qui facit opus Do-

ceitfully, and cursed be he that keepeth back his sword from blood.

mini fraudulenter et maledictus qui prohibet gladium suum

48:11 Moab hath been at ease from his youth, and he hath a sanguine 48:11 fertilis fuit Moab ab adulescentia sua et

settled on his lees, and hath not been emptied from vessel to vessel, neither hath he gone into captivity: therefore his

taste remained in him, and his scent is not changed. requieuit in fecibus suis nec transfusus est de uase in uas et in transmigrationem non abiit idcirco permansit gustus eius

Prophetia Ieremiae

in eo et odor eius non est inmutatus 48:12 propterea ecce 48:12 Therefore, behold, the days come, saith the LORD,

that I will send unto him wanderers, that shall cause him to wander, and shall empty his vessels, and break their

dies ueniunt dicit Dominus et mittam ei ordinatores et stra- bottles. tores laguncularum et sternent eum et uasa eius exhaurient et lagoenas eorum conlident 48:13 et confundetur Moab a 48:13 And Moab shall be ashamed of Chemosh, as the

house of Israel was ashamed of Bethel their confidence.

Chamos sicut confusa est domus Israhel a Bethel in qua habebat fiduciam 48:14 quomodo dicitis fortes sumus et uiri 48:14 How say ye, We are mighty and strong men for the

war?

robusti ad proeliandum 48:15 uastata est Moab et ciuitates 48:15 Moab is spoiled, and gone up out of her cities, and

his chosen young men are gone down to the slaughter, saith the King, whose name is the LORD of hosts.

illius ascenderunt et electi iuuenes eius descenderunt in occisionem ait Rex Dominus exercituum nomen ei 48:16 prope 48:16 The calamity of Moab is near to come, and his af-

fliction hasteth fast.

est interitus Moab ut ueniat et malum eius uelociter adcurret nimis 48:17 consolamini eum omnes qui estis in circuitu 48:17 All ye that are about him, bemoan him; and all ye

that know his name, say, How is the strong staff broken, and the beautiful rod!

eius et uniuersi qui scitis nomen eius dicite quomodo confracta est uirga fortis baculus gloriosus 48:18 descende de 48:18 Thou daughter that dost inhabit Dibon, come down

from thy glory, and sit in thirst; for the spoiler of Moab shall come upon thee, and he shall destroy thy strong

gloria et sede in siti habitatio filiae Dibon quoniam uasta- holds. tor Moab ascendet ad te dissipabit munitiones tuas 48:19 48:19 O inhabitant of Aroer, stand by the way, and espy;

ask him that fleeth, and her that escapeth, and say, What is done?

in uia sta et prospice habitatio Aroer interroga fugientem et eum qui euasit dic quid accidit 48:20 confusus est Moab 48:20 Moab is confounded; for it is broken down: howl

and cry; tell ye it in Arnon, that Moab is spoiled,

quoniam uictus est ululate et clamate adnuntiate in Arnon quoniam uastata est Moab 48:21 et iudicium uenit ad ter- 48:21 And judgment is come upon the plain country; upon

Holon, and upon Jahazah, and upon Mephaath,

ram campestrem super Helon et super Iaesa et super Mefath 48:22 et super Dibon et super Nabo et super domum 48:22 And upon Dibon, and upon Nebo, and upon Bethdi-

blathaim,

Deblathaim 48:23 et super Cariathaim et super Bethgamul 48:23 And upon Kiriathaim, and upon Bethgamul, and

upon Bethmeon,

et super Bethmaon 48:24 et super Carioth et super Bosra et 48:24 And upon Kerioth, and upon Bozrah, and upon all

the cities of the land of Moab, far or near.

super omnes ciuitates terrae Moab quae longe et quae prope sunt 48:25 abscisum est cornu Moab et brachium eius con- 48:25 The horn of Moab is cut off, and his arm is broken,

saith the LORD.

tritum est ait Dominus 48:26 inebriate eum quoniam con- 48:26 Make ye him drunken: for he magnified himself

against the LORD: Moab also shall wallow in his vomit, and he also shall be in derision.

tra Dominum erectus est et adlidet manum Moab in uomitu suo et erit in derisum etiam ipse 48:27 fuit enim in deri- 48:27 For was not Israel a derision unto thee? was he

found among thieves? for since thou spakest of him, thou skippedst for joy.

sum tibi Israhel quasi inter fures repperisses eum propter uerba ergo tua quae aduersum illum locutus es captiuus duceris 48:28 relinquite ciuitates et habitate in petra habitato- 48:28 O ye that dwell in Moab, leave the cities, and dwell

in the rock, and be like the dove that maketh her nest in the sides of the hole’s mouth.

res Moab et estote quasi columba nidificans in summo ore foraminis 48:29 audiuimus superbiam Moab superbus est 48:29 We have heard the pride of Moab, (he is exceeding

proud) his loftiness, and his arrogancy, and his pride, and the haughtiness of his heart.

Biblia Sacra Vulgata

ualde sublimitatem eius et arrogantiam et superbiam et alti-

48:30 I know his wrath, saith the LORD; but it shall not tudinem cordis illius 48:30 ego scio ait Dominus iactantiam

be so; his lies shall not so effect it.

eius et quod non sit iuxta eam uirtus eius nec iuxta quod po-

48:31 Therefore will I howl for Moab, and I will cry terat conata sit facere 48:31 ideo super Moab heiulabo et ad

out for all Moab; mine heart shall mourn for the men of Kirheres.

Moab uniuersam clamabo ad uiros muri fictilis lamentantes

48:32 O vine of Sibmah, I will weep for thee with the 48:32 de planctu Iazer plorabo tibi uinea Sobema propagines

weeping of Jazer: thy plants are gone over the sea, they reach even to the sea of Jazer: the spoiler is fallen upon

thy summer fruits and upon thy vintage. tuae transierunt mare usque ad mare Iazer peruenerunt super 48:33 And joy and gladness is taken from the plentiful messem tuam et uindemiam tuam praedo inruit 48:33 ablata

field, and from the land of Moab, and I have caused wine to fail from the winepresses: none shall tread with shout-

ing; their shouting shall be no shouting. est laetitia et exultatio de Carmelo et de terra Moab et uinum

de torcularibus sustuli nequaquam calcator uuae solitum ce-

48:34 From the cry of Heshbon even unto Elealeh, and leuma cantabit 48:34 de clamore Esebon usque Eleale et

even unto Jahaz, have they uttered their voice, from Zoar even unto Horonaim, as an heifer of three years old: for

the waters also of Nimrim shall be desolate. Iaesa dederunt uocem suam a Segor usque ad Oronaim uitula 48:35 Moreover I will cause to cease in Moab, saith the conternante aquae quoque Namrim pessimae erunt 48:35 et

LORD, him that offereth in the high places, and him that burneth incense to his gods.

auferam de Moab ait Dominus offerentem in excelsis et sa-

48:36 Therefore mine heart shall sound for Moab like crificantem diis eius 48:36 propterea cor meum ad Moab

pipes, and mine heart shall sound like pipes for the men of Kirheres: because the riches that he hath gotten are per-

ished. quasi tibiae resonabit et cor meum ad uiros muri fictilis dabit

sonitum tibiarum quia plus fecit quam potuit idcirco perie-

48:37 For every head shall be bald, and every beard runt 48:37 omne enim caput caluitium et omnis barba rasa

clipped: upon all the hands shall be cuttings, and upon the loins sackcloth.

erit in cunctis manibus conligatio et super omne dorsum cili-

48:38 There shall be lamentation generally upon all the cium 48:38 super omnia tecta Moab et in plateis eius omnis

housetops of Moab, and in the streets thereof: for I have broken Moab like a vessel wherein is no pleasure, saith the

LORD. planctus quia contriui Moab sicut uas inutile ait Dominus 48:39 They shall howl, saying, How is it broken down! 48:39 quomodo uicta est et ululauerunt quomodo deiecit cer-

how hath Moab turned the back with shame! so shall Moab be a derision and a dismaying to all them about him.

uicem Moab et confusus est eritque Moab in derisum et in

48:40 For thus saith the LORD; Behold, he shall fly as an exemplum omnibus in circuitu suo 48:40 haec dicit Domi-

eagle, and shall spread his wings over Moab.

nus ecce quasi aquila euolabit et extendet alas suas ad Moab

48:41 Kerioth is taken, and the strong holds are surprised, 48:41 capta est Carioth et munitiones conprehensae sunt et and the mighty men’s hearts in Moab at that day shall be as the heart of a woman in her pangs.

erit cor fortium Moab in die illa sicut cor mulieris partu-

48:42 And Moab shall be destroyed from being a people, rientis 48:42 et cessabit Moab esse populus quoniam con-

because he hath magnified himself against the LORD. 48:43 Fear, and the pit, and the snare, shall be upon thee, tra Dominum gloriatus est 48:43 pauor et fouea et laqueus O inhabitant of Moab, saith the LORD.

48:44 He that fleeth from the fear shall fall into the pit; super te o habitator Moab ait Dominus 48:44 qui fugit a fa-

and he that getteth up out of the pit shall be taken in the snare: for I will bring upon it, even upon Moab, the year

of their visitation, saith the LORD. cie pauoris cadet in foueam et qui conscenderit de fouea ca-

pietur laqueo adducam enim super Moab annum uisitationis

48:45 They that fled stood under the shadow of Heshbon eorum dicit Dominus 48:45 in umbra Esebon steterunt de

because of the force: but a fire shall come forth out of Heshbon, and a flame from the midst of Sihon, and shall devour the corner of Moab, and the crown of the head of the tumultuous ones.

Prophetia Ieremiae

laqueo fugientes quia ignis egressus est de Esebon et flamma de medio Seon et deuorabit partem Moab et uerticem filiorum tumultus 48:46 uae tibi Moab peristi popule Chamos 48:46 Woe be unto thee, O Moab! the people of Chemosh

perisheth: for thy sons are taken captives, and thy daughters captives.

quia conprehensi sunt filii tui et filiae tuae in captiuitatem 48:47 et conuertam captiuitatem Moab in nouissimis diebus 48:47 Yet will I bring again the captivity of Moab in the

latter days, saith the LORD. Thus far is the judgment of Moab.

ait Dominus hucusque iudicia Moab

49:1 ad filios Ammon haec dicit Dominus numquid filii 49:1 Concerning the Ammonites, thus saith the LORD; Hath Israel no sons? hath he no heir? why then doth their king inherit Gad, and his people dwell in his cities?

non sunt Israhel aut heres non est ei cur igitur hereditate possedit Melchom Gad et populus eius in urbibus eius habitauit 49:2 ideo ecce dies ueniunt dicit Dominus et auditum 49:2 Therefore, behold, the days come, saith the LORD,

that I will cause an alarm of war to be heard in Rabbah of the Ammonites; and it shall be a desolate heap, and her daughters shall be burned with fire: then shall Israel be heir unto them that were his heirs, saith the LORD.

faciam super Rabbath filiorum Ammon fremitum proelii et erit in tumulum dissipata filiaeque eius igni succendentur et possidebit Israhel possessores suos dicit Dominus 49:3 ul- 49:3 Howl, O Heshbon, for Ai is spoiled: cry, ye daughters

of Rabbah, gird you with sackcloth; lament, and run to and fro by the hedges; for their king shall go into captivity, and

ula Esebon quoniam uastata est Ahi clamate filiae Rabbath his priests and his princes together. accingite uos ciliciis plangite et circuite per sepes quia Melchom in transmigratione ducetur sacerdotes eius et principes eius simul 49:4 quid gloriaris in uallibus defluxit uallis tua 49:4 Wherefore gloriest thou in the valleys, thy flowing

valley, O backsliding daughter? that trusted in her treasures, saying, Who shall come unto me?

filia delicata quae confidebas in thesauris tuis et dicebas quis

ueniet ad me 49:5 ecce ego inducam super te terrorem ait 49:5 Behold, I will bring a fear upon thee, saith the Lord GOD of hosts, from all those that be about thee; and ye shall be driven out every man right forth; and none shall

Dominus Deus exercituum ab omnibus qui sunt in circuitu gather up him that wandereth. tuo et dispergemini singuli a conspectu uestro nec erit qui congreget fugientem 49:6 et post haec reuerti faciam cap- 49:6 And afterward I will bring again the captivity of the

children of Ammon, saith the LORD.

tiuos filiorum Ammon ait Dominus 49:7 ad Idumeam haec 49:7 Concerning Edom, thus saith the LORD of hosts; Is wisdom no more in Teman? is counsel perished from the prudent? is their wisdom vanished?

dicit Dominus exercituum numquid non est ultra sapientia in Theman periit consilium a filiis inutilis facta est sapientia eorum 49:8 fugite terga uertite descendite in uoragine habi- 49:8 Flee ye, turn back, dwell deep, O inhabitants of

Dedan; for I will bring the calamity of Esau upon him, the time that I will visit him.

tatores Dedan quoniam perditionem Esau adduxi super eum tempus uisitationis eius 49:9 si uindemiatores uenissent su- 49:9 If grapegatherers come to thee, would they not leave

some gleaning grapes? if thieves by night, they will destroy till they have enough.

per te non reliquissent racemum si fures in nocte rapuissent quod sufficeret sibi 49:10 ego uero discoperui Esau reuelaui 49:10 But I have made Esau bare, I have uncovered his

secret places, and he shall not be able to hide himself: his seed is spoiled, and his brethren, and his neighbours, and

abscondita eius et celari non poterit uastatum est semen eius he is not. et fratres eius et uicini eius et non erit 49:11 relinque pupil- 49:11 Leave thy fatherless children, I will preserve them

alive; and let thy widows trust in me.

los tuos ego eos faciam uiuere et uiduae tuae in me sperabunt

Biblia Sacra Vulgata

49:12 For thus saith the LORD; Behold, they whose judg- 49:12 quia haec dicit Dominus ecce quibus non erat iudicium

ment was not to drink of the cup have assuredly drunken; and art thou he that shall altogether go unpunished? thou

shalt not go unpunished, but thou shalt surely drink of it. ut biberent calicem bibentes bibent et tu quasi innocens re-49:13 For I have sworn by myself, saith the LORD, that linqueris non eris innocens sed bibens bibes 49:13 quia per

Bozrah shall become a desolation, a reproach, a waste, and a curse; and all the cities thereof shall be perpetual wastes.

memet ipsum iuraui dicit Dominus quod in solitudinem et in obprobrium et in desertum et in maledictionem erit Bosra et

49:14 I have heard a rumour from the LORD, and an am- omnes ciuitates eius erunt in solitudines sempiternas 49:14

bassador is sent unto the heathen, saying, Gather ye together, and come against her, and rise up to the battle.

auditum audiui a Domino et legatus ad gentes missus est congregamini et uenite contra eam et consurgamus in pro-

49:15 For, lo, I will make thee small among the heathen, elium 49:15 ecce enim paruulum dedi te in gentibus conand despised among men. 49:16 Thy terribleness hath deceived thee, and the pride of temptibilem inter homines 49:16 arrogantia tua decepit te thine heart, O thou that dwellest in the clefts of the rock, that holdest the height of the hill: though thou shouldest make thy nest as high as the eagle, I will bring thee down from thence, saith the LORD.

et superbia cordis tui qui habitas in cauernis petrae et adprehendere niteris altitudinem collis cum exaltaueris quasi

49:17 Also Edom shall be a desolation: every one that aquila nidum tuum inde detraham te dicit Dominus 49:17

goeth by it shall be astonished, and shall hiss at all the plagues thereof.

et erit Idumea deserta omnis qui transibit per eam stupebit

49:18 As in the overthrow of Sodom and Gomorrah and et sibilabit super omnes plagas eius 49:18 sicuti subuersa

the neighbour cities thereof, saith the LORD, no man shall abide there, neither shall a son of man dwell in it.

est Sodoma et Gomorra et uicinae eius ait Dominus non ha-

49:19 Behold, he shall come up like a lion from the bitabit ibi uir et non incolet eam filius hominis 49:19 ecce

swelling of Jordan against the habitation of the strong: but I will suddenly make him run away from her: and who is a chosen man, that I may appoint over her? for who is like me? and who will appoint me the time? and who is that shepherd that will stand before me?

quasi leo ascendet de superbia Iordanis ad pulchritudinem robustam quia subito currere eum faciam ad illam et quis erit electus quem praeponam ei quis enim similis mei et quis sustinebit me et quis est iste pastor qui resistat uultui meo

49:20 Therefore hear the counsel of the LORD, that he 49:20 propterea audite consilium Domini quod iniit de Edom

hath taken against Edom; and his purposes, that he hath purposed against the inhabitants of Teman: Surely the least of the flock shall draw them out: surely he shall make their habitations desolate with them.

et cogitationes eius quas cogitauit de habitatoribus Theman si non deiecerint eos paruuli gregis nisi dissipauerint cum eis

49:21 The earth is moved at the noise of their fall, at the habitaculum eorum 49:21 a uoce ruinae eorum commota est cry the noise thereof was heard in the Red sea.

49:22 Behold, he shall come up and fly as the eagle, and terra clamor in mari Rubro auditus est uocis eius 49:22 ecce

spread his wings over Bozrah: and at that day shall the heart of the mighty men of Edom be as the heart of a

woman in her pangs. quasi aquila ascendet et euolabit et expandet alas suas super

Bosram et erit cor fortium Idumeae in die illa quasi cor mu-

49:23 Concerning Damascus. Hamath is confounded, and lieris parturientis 49:23 ad Damascum confusa est Emath et

Arpad: for they have heard evil tidings: they are fainthearted; there is sorrow on the sea; it cannot be quiet.

Arfad quia auditum pessimum audierunt turbati sunt in mari

49:24 Damascus is waxed feeble, and turneth herself to sollicitudine quiescere non potuit 49:24 dissoluta est Da-

flee, and fear hath seized on her: anguish and sorrows have taken her, as a woman in travail.

mascus uersa in fugam tremor adprehendit eam angustia et

49:25 How is the city of praise not left, the city of my joy! dolores tenuerunt eam quasi parturientem 49:25 quomodo

Prophetia Ieremiae

dereliquerunt ciuitatem laudabilem urbem laetitiae 49:26 49:26 Therefore her young men shall fall in her streets,

and all the men of war shall be cut off in that day, saith the LORD of hosts.

ideo cadent iuuenes eius in plateis eius et omnes uiri proelii conticescent in die illa ait Dominus exercituum 49:27 49:27 And I will kindle a fire in the wall of Damascus, and

it shall consume the palaces of Benhadad.

et succendam ignem in muro Damasci et deuorabit moenia Benadad 49:28 ad Cedar et ad regna Asor quae percussit 49:28 Concerning Kedar, and concerning the kingdoms

of Hazor, which Nebuchadrezzar king of Babylon shall smite, thus saith the LORD; Arise ye, go up to Kedar, and

Nabuchodonosor rex Babylonis haec dicit Dominus surgite spoil the men of the east. ascendite ad Cedar et uastate filios orientis 49:29 taberna- 49:29 Their tents and their flocks shall they take away:

they shall take to themselves their curtains, and all their vessels, and their camels; and they shall cry unto them,

cula eorum et greges eorum capient pelles eorum et omnia Fear is on every side. uasa eorum et camelos eorum tollent sibi et uocabunt super eos formidinem in circuitu 49:30 fugite abite uehemen- 49:30 Flee, get you far off, dwell deep, O ye inhabitants of

Hazor, saith the LORD; for Nebuchadrezzar king of Babylon hath taken counsel against you, and hath conceived a

ter in uoraginibus sedete qui habitatis Asor ait Dominus in- purpose against you. iit enim contra uos Nabuchodonosor rex Babylonis consilium et cogitauit aduersum uos cogitationes 49:31 consur- 49:31 Arise, get you up unto the wealthy nation, that

dwelleth without care, saith the LORD, which have neither gates nor bars, which dwell alone.

gite et ascendite ad gentem quietam et habitantem confidenter ait Dominus non ostia non uectes ei soli habitant 49:32 49:32 And their camels shall be a booty, and the multi-

tude of their cattle a spoil: and I will scatter into all winds them that are in the utmost corners; and I will bring their

et erunt cameli eorum in direptionem et multitudo iumen- calamity from all sides thereof, saith the LORD. torum in praedam et dispergam eos in omnem uentum qui sunt adtonsi in comam et ex omni confinio eorum adducam interitum super eos ait Dominus 49:33 et erit Asor in habi- 49:33 And Hazor shall be a dwelling for dragons, and a

desolation for ever: there shall no man abide there, nor any son of man dwell in it.

taculum draconum deserta usque in aeternum non manebit ibi uir nec incolet eam filius hominis 49:34 quod factum est 49:34 The word of the LORD that came to Jeremiah the

prophet against Elam in the beginning of the reign of Zedekiah king of Judah, saying,

uerbum Domini ad Hieremiam prophetam aduersus Aelam in principio regni Sedeciae regis Iuda dicens 49:35 haec di- 49:35 Thus saith the LORD of hosts; Behold, I will break

the bow of Elam, the chief of their might.

cit Dominus exercituum ecce ego confringam arcum Aelam summam fortitudinem eorum 49:36 et inducam super Ae- 49:36 And upon Elam will I bring the four winds from

the four quarters of heaven, and will scatter them toward all those winds; and there shall be no nation whither the

lam quattuor uentos a quattuor plagis caeli et uentilabo eos outcasts of Elam shall not come. in omnes uentos istos et non erit gens ad quam non perueniant profugi Aelam 49:37 et pauere faciam Aelam coram 49:37 For I will cause Elam to be dismayed before their

enemies, and before them that seek their life: and I will bring evil upon them, even my fierce anger, saith the LORD; and I will send the sword after them, till I have consumed them:

inimicis suis et in conspectu quaerentium animam eorum et adducam super eos malum iram furoris mei dicit Dominus et emittam post eos gladium donec consumam eos 49:38 et 49:38 And I will set my throne in Elam, and will destroy

from thence the king and the princes, saith the LORD.

ponam solium meum in Aelam et perdam inde reges et principes ait Dominus 49:39 in nouissimis autem diebus reuerti 49:39 But it shall come to pass in the latter days, that I will

bring again the captivity of Elam, saith the LORD.

Biblia Sacra Vulgata

faciam captiuos Aelam dicit Dominus

50:1 The word that the LORD spake against Babylon and 50:1 uerbum quod locutus est Dominus de Babylone et de

against the land of the Chaldeans by Jeremiah the prophet. 50:2 Declare ye among the nations, and publish, and set terra Chaldeorum in manu Hieremiae prophetae 50:2 adup a standard; publish, and conceal not: say, Babylon is taken, Bel is confounded, Merodach is broken in pieces;

her idols are confounded, her images are broken in pieces. nuntiate in gentibus et auditum facite leuate signum praedicate et nolite celare dicite capta est Babylon confusus est Bel uictus est Marodach confusa sunt sculptilia eius superata

50:3 For out of the north there cometh up a nation against sunt idola eorum 50:3 quoniam ascendit contra eam gens her, which shall make her land desolate, and none shall dwell therein: they shall remove, they shall depart, both

man and beast. ab aquilone quae ponet terram eius in solitudinem et non

erit qui habitet in ea ab homine usque ad pecus et moti sunt

50:4 In those days, and in that time, saith the LORD, the et abierunt 50:4 in diebus illis et in tempore illo ait Do-

children of Israel shall come, they and the children of Judah together, going and weeping: they shall go, and seek

the LORD their God. minus uenient filii Israhel ipsi et filii Iuda simul ambulan-

tes et flentes properabunt et Dominum Deum suum quae-

50:5 They shall ask the way to Zion with their faces thith- rent 50:5 in Sion interrogabunt uiam huc facies eorum ueerward, saying, Come, and let us join ourselves to the LORD in a perpetual covenant that shall not be forgotten.

nient et adponentur ad Dominum foedere sempiterno quod

50:6 My people hath been lost sheep: their shepherds have nulla obliuione delebitur 50:6 grex perditus factus est popu-

caused them to go astray, they have turned them away on the mountains: they have gone from mountain to hill, they

have forgotten their restingplace. lus meus pastores eorum seduxerunt eos feceruntque uagari

in montibus de monte in collem transierunt obliti sunt cubi-

50:7 All that found them have devoured them: and their lis sui 50:7 omnes qui inuenerunt comederunt eos et hostes

adversaries said, We offend not, because they have sinned against the LORD, the habitation of justice, even the

LORD, the hope of their fathers. eorum dixerunt non peccauimus pro eo quod peccauerunt

Domino decori iustitiae et expectationi patrum eorum Do-

50:8 Remove out of the midst of Babylon, and go forth out mino 50:8 recedite de medio Babylonis et de terra Chalde-

of the land of the Chaldeans, and be as the he goats before the flocks.

orum egredimini et estote quasi hedi ante greges 50:9 quo-

50:9 For, lo, I will raise and cause to come up against Babylon an assembly of great nations from the north country: and they shall set themselves in array against her; from thence she shall be taken: their arrows shall be as of a mighty expert man; none shall return in vain.

niam ecce ego suscito et adducam in Babylonem congregationem gentium magnarum de terra aquilonis et praeparabuntur aduersum eam et inde capietur sagitta eius quasi uiri

50:10 And Chaldea shall be a spoil: all that spoil her shall fortis interfectoris non reuertetur uacua 50:10 et erit Chal-

be satisfied, saith the LORD.

dea in praedam omnes uastantes eam replebuntur ait Domi-

50:11 Because ye were glad, because ye rejoiced, O ye nus 50:11 quoniam exultatis et magna loquimini diripientes

destroyers of mine heritage, because ye are grown fat as the heifer at grass, and bellow as bulls;

hereditatem meam quoniam effusi estis sicut uitulus super

50:12 Your mother shall be sore confounded; she that bare herbam et mugistis ut tauri 50:12 confusa est mater uestra

you shall be ashamed: behold, the hindermost of the nations shall be a wilderness, a dry land, and a desert.

nimis et adaequata pulueri quae genuit uos ecce nouissima

50:13 Because of the wrath of the LORD it shall not be erit in gentibus deserta inuia et arens 50:13 ab ira Domini

inhabited, but it shall be wholly desolate: every one that goeth by Babylon shall be astonished, and hiss at all her

plagues. non habitabitur sed rediietur tota in solitudinem omnis qui

Prophetia Ieremiae

transit per Babylonem stupebit et sibilabit super uniuersis plagis eius 50:14 praeparamini contra Babylonem per cir- 50:14 Put yourselves in array against Babylon round

about: all ye that bend the bow, shoot at her, spare no arrows: for she hath sinned against the LORD.

cuitum omnes qui intenditis arcum debellate eam non parcatis iaculis quia Domino peccauit 50:15 clamate aduersus 50:15 Shout against her round about: she hath given

her hand: her foundations are fallen, her walls are thrown down: for it is the vengeance of the LORD: take

eam ubique dedit manum ceciderunt fundamenta eius de- vengeance upon her; as she hath done, do unto her. structi sunt muri eius quoniam ultio Domini est ultionem accipite de ea sicut fecit facite ei 50:16 disperdite satorem de 50:16 Cut off the sower from Babylon, and him that han-

dleth the sickle in the time of harvest: for fear of the oppressing sword they shall turn every one to his people, and

Babylone et tenentem falcem in tempore messis a facie gla- they shall flee every one to his own land. dii columbae unusquisque ad populum suum conuertetur et

singuli ad terram suam fugient 50:17 grex dispersus Isra- 50:17 Israel is a scattered sheep; the lions have driven him away: first the king of Assyria hath devoured him; and last this Nebuchadrezzar king of Babylon hath broken his

hel leones eiecerunt eum primus comedit eum rex Assur iste bones. nouissimus exossauit eum Nabuchodonosor rex Babylonis 50:18 propterea haec dicit Dominus exercituum Deus Isra- 50:18 Therefore thus saith the LORD of hosts, the God of

Israel; Behold, I will punish the king of Babylon and his land, as I have punished the king of Assyria.

hel ecce ego uisitabo regem Babylonis et terram eius sicut uisitaui regem Assur 50:19 et reducam Israhel ad habitacu- 50:19 And I will bring Israel again to his habitation, and

he shall feed on Carmel and Bashan, and his soul shall be satisfied upon mount Ephraim and Gilead.

lum suum et pascetur Carmelum et Basan et in monte Ephraim et Galaad saturabitur anima eius 50:20 in diebus illis 50:20 In those days, and in that time, saith the LORD, the

iniquity of Israel shall be sought for, and there shall be none; and the sins of Judah, and they shall not be found:

et in tempore illo ait Dominus quaeretur iniquitas Israhel et for I will pardon them whom I reserve. non erit et peccatum Iuda et non inuenietur quoniam propitius ero eis quos reliquero 50:21 super terram dominantium 50:21 Go up against the land of Merathaim, even against

it, and against the inhabitants of Pekod: waste and utterly destroy after them, saith the LORD, and do according to

ascende et super habitatores eius uisita dissipa et interfice all that I have commanded thee. quae post eos sunt ait Dominus et fac iuxta omnia quae praecepi tibi 50:22 uox belli in terra et contritio magna 50:23 50:22 A sound of battle is in the land, and of great destruc-

tion.

50:23 How is the hammer of the whole earth cut asunder and broken! how is Babylon become a desolation among the nations!

quomodo confractus est et contritus est malleus uniuersae terrae quomodo uersa est in desertum Babylon in gentibus 50:24 inlaqueaui te et capta es Babylon et nesciebas inuenta 50:24 I have laid a snare for thee, and thou art also taken,

O Babylon, and thou wast not aware: thou art found, and also caught, because thou hast striven against the LORD.

es et adprehensa quoniam Dominum prouocasti 50:25 ape-

50:25 The LORD hath opened his armoury, and hath brought forth the weapons of his indignation: for this is the work of the Lord GOD of hosts in the land of the Chaldeans.

ruit Dominus thesaurum suum et protulit uasa irae suae quoniam opus est Domino Deo exercituum in terra Chaldeorum 50:26 uenite ad eam ab extremis finibus aperite ut exeant qui 50:26 Come against her from the utmost border, open her

storehouses: cast her up as heaps, and destroy her utterly: let nothing of her be left.

conculcent eam tollite de uia lapides et rediiite in aceruos et interficite eam nec sit quicquam reliquum 50:27 dissi- 50:27 Slay all her bullocks; let them go down to the

slaughter: woe unto them! for their day is come, the time of their visitation.

pate uniuersos fortes eius descendant in occisionem uae eis

Biblia Sacra Vulgata

50:28 The voice of them that flee and escape out of the quia uenit dies eorum tempus uisitationis eorum 50:28 uox

land of Babylon, to declare in Zion the vengeance of the LORD our God, the vengeance of his temple.

fugientium et eorum qui euaserunt de terra Babylonis ut adnuntient in Sion ultionem Domini Dei nostri ultionem templi

50:29 Call together the archers against Babylon: all ye eius 50:29 adnuntiate in Babylonem plurimis omnibus qui

that bend the bow, camp against it round about; let none thereof escape: recompense her according to her work; according to all that she hath done, do unto her: for she hath been proud against the LORD, against the Holy One of Israel.

tendunt arcum consistite aduersum eam per gyrum et nullus euadat reddite ei secundum opus suum iuxta omnia quae fecit facite illi quia contra Dominum erecta est aduersum

50:30 Therefore shall her young men fall in the streets, Sanctum Israhel 50:30 idcirco cadent iuuenes eius in plateis

and all her men of war shall be cut off in that day, saith the LORD.

eius et omnes uiri bellatores eius conticescent in die illa ait

50:31 Behold, I am against thee, O thou most proud, saith Dominus 50:31 ecce ego ad te superbe dicit Dominus Dethe Lord GOD of hosts: for thy day is come, the time that I will visit thee.

us exercituum quia uenit dies tuus tempus uisitationis tuae

50:32 And the most proud shall stumble and fall, and none 50:32 et cadet superbus et corruet et non erit qui suscitet eum

shall raise him up: and I will kindle a fire in his cities, and it shall devour all round about him.

et succendam ignem in urbibus eius et deuorabit omnia in

50:33 Thus saith the LORD of hosts; The children of Israel circuitu eius 50:33 haec dicit Dominus exercituum calum-

and the children of Judah were oppressed together: and all that took them captives held them fast; they refused to let

them go. niam sustinent filii Israhel et filii Iuda simul omnes qui ce-50:34 Their Redeemer is strong; the LORD of hosts is his perunt eos tenent nolunt dimittere eos 50:34 redeeptor eo-

name: he shall throughly plead their cause, that he may give rest to the land, and disquiet the inhabitants of Baby-

lon. rum Fortis Dominus exercituum nomen eius iudicio defen-

det causam eorum ut exterreat terram et commoueat habita-

50:35 A sword is upon the Chaldeans, saith the LORD, tores Babylonis 50:35 gladius ad Chaldeos ait Dominus et

and upon the inhabitants of Babylon, and upon her princes, and upon her wise men.

ad habitatores Babylonis et ad principes et ad sapientes eius

50:36 A sword is upon the liars; and they shall dote: a 50:36 gladius ad diuinos eius qui stulti erunt gladius ad fortes

sword is upon her mighty men; and they shall be dismayed.

illius qui timebunt 50:37 gladius ad equos eius et ad currus

50:37 A sword is upon their horses, and upon their chariots, and upon all the mingled people that are in the midst of her; and they shall become as women: a sword is upon her treasures; and they shall be robbed.

eius et ad omne uulgus quod est in medio eius et erunt quasi

50:38 A drought is upon her waters; and they shall be dried mulieres gladius ad thesauros eius qui diripientur 50:38 sic-

up: for it is the land of graven images, and they are mad upon their idols.

citas super aquas eius erit et arescent quia terra sculptilium

50:39 Therefore the wild beasts of the desert with the wild est et in portentis gloriantur 50:39 propterea habitabunt dra-

beasts of the islands shall dwell there, and the owls shall dwell therein: and it shall be no more inhabited for ever;

neither shall it be dwelt in from generation to generation. cones cum fatuis ficariis et habitabunt in ea strutiones et non

habitabitur ultra usque ad sempiternum nec extruetur usque

50:40 As God overthrew Sodom and Gomorrah and the ad generationem et generationem 50:40 sicut subuertit De-

neighbour cities thereof, saith the LORD; so shall no man abide there, neither shall any son of man dwell therein.

us Sodomam et Gomorram et uicinas eius ait Dominus non

50:41 Behold, a people shall come from the north, and a habitabit ibi uir nec incolet eam filius hominis 50:41 ecce

great nation, and many kings shall be raised up from the coasts of the earth.

populus uenit ab aquilone et gens magna et reges multi con-

50:42 They shall hold the bow and the lance: they are surgent a finibus terrae 50:42 arcum et scutum adprehen-

cruel, and will not shew mercy: their voice shall roar like the sea, and they shall ride upon horses, every one put in array, like a man to the battle, against thee, O daughter of Babylon.

Prophetia Ieremiae

dent crudeles sunt et inmisericordes uox eorum quasi mare sonabit et super equos ascendent sicut uir paratus ad proelium contra te filia Babylon 50:43 audiuit rex Babylonis 50:43 The king of Babylon hath heard the report of them,

and his hands waxed feeble: anguish took hold of him, and pangs as of a woman in travail.

famam eorum et dissolutae sunt manus eius angustia adprehendit eum dolor quasi parturientem 50:44 ecce quasi leo 50:44 Behold, he shall come up like a lion from the

swelling of Jordan unto the habitation of the strong: but I will make them suddenly run away from her: and who is a chosen man, that I may appoint over her? for who is like me? and who will appoint me the time? and who is that shepherd that will stand before me?

ascendet de superbia Iordanis ad pulchritudinem robustam quia subito currere eum faciam ad illam et quis erit electus quem praeponam ei quis enim similis mei et quis sustinebit me et quis est iste pastor qui resistat uultui meo 50:45 50:45 Therefore hear ye the counsel of the LORD, that

he hath taken against Babylon; and his purposes, that he hath purposed against the land of the Chaldeans: Surely the least of the flock shall draw them out: surely he shall make their habitation desolate with them.

propterea audite consilium Domini quod mente concepit aduersum Babylonem et cogitationes eius quas cogitauit super terram Chaldeorum nisi detraxerint eos paruuli gregum nisi dissipatum fuerit cum ipsis habitaculum eorum 50:46 a 50:46 At the noise of the taking of Babylon the earth is

moved, and the cry is heard among the nations.

uoce captiuitatis Babylonis commota est terra et clamor inter gentes auditus est

51:1 haec dicit Dominus ecce ego suscitabo super Babylo- 51:1 Thus saith the LORD; Behold, I will raise up against

Babylon, and against them that dwell in the midst of them that rise up against me, a destroying wind;

nem et super habitatores eius qui cor suum leuauerunt contra me quasi uentum pestilentem 51:2 et mittam in Babylonem 51:2 And will send unto Babylon fanners, that shall fan

her, and shall empty her land: for in the day of trouble they shall be against her round about.

uentilatores et uentilabunt eam et demolientur terram eius quoniam uenerunt super eam undique in die adflictionis eius

51:3 non tendat qui tendit arcum suum et non ascendat lo- 51:3 Against him that bendeth let the archer bend his bow, and against him that lifteth himself up in his brigandine: and spare ye not her young men; destroy ye utterly all her

ricatus nolite parcere iuuenibus eius interficite omnem mi- host. litiam eius 51:4 et cadent interfecti in terra Chaldeorum et 51:4 Thus the slain shall fall in the land of the Chaldeans,

and they that are thrust through in her streets.

uulnerati in regionibus eius 51:5 quoniam non fuit uiduatus 51:5 For Israel hath not been forsaken, nor Judah of his

God, of the LORD of hosts; though their land was filled with sin against the Holy One of Israel.

Israhel et Iuda a Deo suo Domino exercituum terra autem eorum repleta est delicto a Sancto Israhel 51:6 fugite de 51:6 Flee out of the midst of Babylon, and deliver every

man his soul: be not cut off in her iniquity; for this is the time of the LORD’s vengeance; he will render unto her a

medio Babylonis et saluet unusquisque animam suam no- recompence. lite tacere super iniquitatem eius quoniam tempus ultionis est Domino uicissitudinem ipse retribuet ei 51:7 calix au- 51:7 Babylon hath been a golden cup in the LORD’s hand,

that made all the earth drunken: the nations have drunken of her wine; therefore the nations are mad.

reus Babylon in manu Domini inebrians omnem terram de uino eius biberunt gentes et ideo commotae sunt 51:8 su- 51:8 Babylon is suddenly fallen and destroyed: howl for

her; take balm for her pain, if so be she may be healed.

bito cecidit Babylon et contrita est ululate super eam tollite resinam ad dolorem eius si forte sanetur 51:9 curaui- 51:9 We would have healed Babylon, but she is not healed:

forsake her, and let us go every one into his own country: for her judgment reacheth unto heaven, and is lifted up even to the skies.

Biblia Sacra Vulgata

mus Babylonem et non est sanata derelinquamus eam et eamus unusquisque in terram suam quoniam peruenit usque ad

51:10 The LORD hath brought forth our righteousness: caelos iudicium eius et eleuatum est usque ad nubes 51:10

come, and let us declare in Zion the work of the LORD our God.

protulit Dominus iustitias nostras uenite et narremus in Sion

51:11 Make bright the arrows; gather the shields: the opus Domini Dei nostri 51:11 acuite sagittas implete fare-

LORD hath raised up the spirit of the kings of the Medes: for his device is against Babylon, to destroy it; because it is the vengeance of the LORD, the vengeance of his temple.

tras suscitauit Dominus spiritum regum Medorum et contra Babylonem mens eius ut perdat eam quoniam ultio Domini

51:12 Set up the standard upon the walls of Babylon, make est ultio templi sui 51:12 super muros Babylonis leuate si-

the watch strong, set up the watchmen, prepare the ambushes: for the LORD hath both devised and done that

which he spake against the inhabitants of Babylon. gnum augete custodiam leuate custodes praeparate insidias

quia cogitauit Dominus et fecit quaecumque locutus est con-

51:13 O thou that dwellest upon many waters, abundant tra habitatores Babylonis 51:13 quae habitas super aquas

in treasures, thine end is come, and the measure of thy covetousness.

multas locuples in thesauris uenit finis tuus pedalis praeci-

51:14 The LORD of hosts hath sworn by himself, saying, sionis tuae 51:14 iurauit Dominus exercituum per animam

Surely I will fill thee with men, as with caterpillers; and they shall lift up a shout against thee.

suam quoniam replebo te hominibus quasi brucho et super te

51:15 He hath made the earth by his power, he hath estab- celeuma cantabitur 51:15 qui fecit terram in fortitudine sua

lished the world by his wisdom, and hath stretched out the heaven by his understanding.

praeparauit orbem in sapientia sua et prudentia sua extendit

51:16 When he uttereth his voice, there is a multitude of caelos 51:16 dante eo uocem multiplicantur aquae in caelo

waters in the heavens; and he causeth the vapours to ascend from the ends of the earth: he maketh lightnings with

rain, and bringeth forth the wind out of his treasures. qui leuat nubes ab extremo terrae fulgura in pluuiam fecit 51:17 Every man is brutish by his knowledge; every et produxit uentum de thesauris suis 51:17 stultus factus est

founder is confounded by the graven image: for his molten image is falsehood, and there is no breath in them.

omnis homo ab scientia confusus est omnis conflator in scul-

51:18 They are vanity, the work of errors: in the time of ptili quia mendax conflatio eius nec est spiritus in eis 51:18

their visitation they shall perish.

uana sunt opera et risu digna in tempore uisitationis suae per-

51:19 The portion of Jacob is not like them; for he is the ibunt 51:19 non sicut haec pars Iacob quia qui fecit omnia

former of all things: and Israel is the rod of his inheritance: the LORD of hosts is his name.

ipse est et Israhel sceptrum hereditatis eius Dominus exer-

51:20 Thou art my battle axe and weapons of war: for with cituum nomen eius 51:20 conlidis tu mihi uasa belli et ego

thee will I break in pieces the nations, and with thee will I destroy kingdoms;

conlidam in te gentes et disperdam in te regna 51:21 et con-

51:21 And with thee will I break in pieces the horse and his rider; and with thee will I break in pieces the chariot and his rider;

lidam in te equum et equitem eius et conlidam in te currum et

51:22 With thee also will I break in pieces man and ascensorem eius 51:22 et conlidam in te uirum et mulierem

woman; and with thee will I break in pieces old and young; and with thee will I break in pieces the young man and the

maid; et conlidam in te senem et puerum et conlidam in te iuuenem 51:23 I will also break in pieces with thee the shepherd et uirginem 51:23 et conlidam in te pastorem et gregem eius

and his flock; and with thee will I break in pieces the husbandman and his yoke of oxen; and with thee will I break

in pieces captains and rulers. et conlidam in te agricolam et iugales eius et conlidam in 51:24 And I will render unto Babylon and to all the inhab- te duces et magistratus 51:24 et reddam Babyloni et cunc-

itants of Chaldea all their evil that they have done in Zion in your sight, saith the LORD.

tis habitatoribus Chaldeae omne malum suum quod fecerunt

Prophetia Ieremiae

in Sion in oculis uestris ait Dominus 51:25 ecce ego ad te 51:25 Behold, I am against thee, O destroying mountain,

saith the LORD, which destroyest all the earth: and I will stretch out mine hand upon thee, and roll thee down from

mons pestifer ait Dominus qui corrumpis uniuersam terram the rocks, and will make thee a burnt mountain. et extendam manum meam super te et euoluam te de petris et dabo te in montem conbustionis 51:26 et non tollent de 51:26 And they shall not take of thee a stone for a corner,

nor a stone for foundations; but thou shalt be desolate for ever, saith the LORD.

te lapidem in angulum et lapidem in fundamenta sed perditus in aeternum eris ait Dominus 51:27 leuate signum in 51:27 Set ye up a standard in the land, blow the trumpet

among the nations, prepare the nations against her, call together against her the kingdoms of Ararat, Minni, and Ashchenaz; appoint a captain against her; cause the horses to come up as the rough caterpillers.

terra clangite bucina in gentibus sanctificate super eam gentes adnuntiate contra illam regibus Ararat Menni et Aschenez numerate contra eam Thapsar adducite equum quasi bruchum aculeatum 51:28 sanctificate contra eam gentes reges 51:28 Prepare against her the nations with the kings of the

Medes, the captains thereof, and all the rulers thereof, and all the land of his dominion.

Mediae duces eius et uniuersos magistratus eius cunctamque terram potestatis eius 51:29 et commouebitur terra et 51:29 And the land shall tremble and sorrow: for every

purpose of the LORD shall be performed against Babylon, to make the land of Babylon a desolation without an

turbabitur quia euigilauit contra Babylonem cogitatio Do- inhabitant. mini ut ponat terram Babylonis desertam et inhabitabilem 51:30 cessauerunt fortes Babylonis a proelio habitauerunt in 51:30 The mighty men of Babylon have forborn to fight,

they have remained in their holds: their might hath failed; they became as women: they have burned her dwelling-

praesidiis deuoratum est robur eorum et facti sunt quasi mu- places; her bars are broken. lieres incensa sunt tabernacula eius contriti sunt uectes eius 51:31 currens obuiam currenti ueniet et nuntius obuius nun- 51:31 One post shall run to meet another, and one messen-

ger to meet another, to shew the king of Babylon that his city is taken at one end,

tianti ut adnuntiet regi Babylonis quia capta est ciuitas eius a summo usque ad summum 51:32 et uada praeoccupata sunt 51:32 And that the passages are stopped, and the reeds

they have burned with fire, and the men of war are affrighted.

et paludes incensae sunt igni et uiri bellatores conturbati sunt 51:33 quia haec dicit Dominus exercituum Deus Israhel fi- 51:33 For thus saith the LORD of hosts, the God of Israel;

The daughter of Babylon is like a threshingfloor, it is time to thresh her: yet a little while, and the time of her harvest

lia Babylon quasi area tempus triturae eius adhuc modicum shall come. et ueniet tempus messionis eius 51:34 comedit me deuora- 51:34 Nebuchadrezzar the king of Babylon hath devoured

me, he hath crushed me, he hath made me an empty vessel, he hath swallowed me up like a dragon, he hath filled his

uit me Nabuchodonosor rex Babylonis reddidit me quasi uas belly with my delicates, he hath cast me out. inane absorbuit me sicut draco repleuit uentrem suum teneritudine mea et eiecit me 51:35 iniquitas aduersum me et caro 51:35 The violence done to me and to my flesh be upon

Babylon, shall the inhabitant of Zion say; and my blood upon the inhabitants of Chaldea, shall Jerusalem say.

mea super Babylonem dicit habitatio Sion et sanguis meus super habitatores Chaldeae dicit Hierusalem 51:36 propte- 51:36 Therefore thus saith the LORD; Behold, I will plead

thy cause, and take vengeance for thee; and I will dry up her sea, and make her springs dry.

rea haec dicit Dominus ecce ego iudicabo causam tuam et ulciscar ultionem tuam et desertum faciam mare eius et siccabo uenam eius 51:37 et erit Babylon in tumulos habitatio 51:37 And Babylon shall become heaps, a dwellingplace

for dragons, an astonishment, and an hissing, without an inhabitant.

draconum stupor et sibilus eo quod non sit habitator 51:38

51:38 They shall roar together like lions: they shall yell as lions’ whelps.

Biblia Sacra Vulgata

simul ut leones rugient excutient comas uelut catuli leonum

51:39 In their heat I will make their feasts, and I will make 51:39 in calore eorum ponam potus eorum et inebriabo eos them drunken, that they may rejoice, and sleep a perpetual sleep, and not wake, saith the LORD.

ut sopiantur et dormiant somnum sempiternum et non con-

51:40 I will bring them down like lambs to the slaughter, surgant dicit Dominus 51:40 deducam eos quasi agnos ad like rams with he goats.

51:41 How is Sheshach taken! and how is the praise of the uictimam quasi arietes cum hedis 51:41 quomodo capta est

whole earth surprised! how is Babylon become an astonishment among the nations!

Sesach et conprehensa est inclita uniuersae terrae quomodo

51:42 The sea is come up upon Babylon: she is covered facta est in stuporem Babylon inter gentes 51:42 ascendit

with the multitude of the waves thereof.

super Babylonem mare multitudine fluctuum eius operta est

51:43 Her cities are a desolation, a dry land, and a wilder- 51:43 factae sunt ciuitates eius in stuporem terra inhabitabi-

ness, a land wherein no man dwelleth, neither doth any son of man pass thereby.

lis et deserta terra in qua nullus habitet nec transeat per eam

51:44 And I will punish Bel in Babylon, and I will bring filius hominis 51:44 et uisitabo super Bel in Babylone et ei-

forth out of his mouth that which he hath swallowed up: and the nations shall not flow together any more unto him:

yea, the wall of Babylon shall fall. ciam quod absorbuerat de ore eius et non confluent ad eum 51:45 My people, go ye out of the midst of her, and deliver ultra gentes siquidem et murus Babylonis corruit 51:45 eg-

ye every man his soul from the fierce anger of the LORD.

redimini de medio eius populus meus ut saluet unusquisque

51:46 And lest your heart faint, and ye fear for the rumour animam suam ab ira furoris Domini 51:46 et ne forte molthat shall be heard in the land; a rumour shall both come one year, and after that in another year shall come a ru-

mour, and violence in the land, ruler against ruler. lescat cor uestrum et timeatis auditum qui audietur in terra

et ueniet in anno auditio et post hunc annum auditio et in-

51:47 Therefore, behold, the days come, that I will do iquitas in terra et dominator super dominatorem 51:47 pro-

judgment upon the graven images of Babylon: and her whole land shall be confounded, and all her slain shall fall

in the midst of her. pterea ecce dies ueniunt et uisitabo super sculptilia Babylo-

nis et omnis terra eius confundetur et uniuersi interfecti eius

51:48 Then the heaven and the earth, and all that is therein, cadent in medio eius 51:48 et laudabunt super Babylonem shall sing for Babylon: for the spoilers shall come unto her from the north, saith the LORD.

caeli et terra et omnia quae in eis sunt quia ab aquilone ue-

51:49 As Babylon hath caused the slain of Israel to fall, so nient ei praedones ait Dominus 51:49 et quomodo fecit Baat Babylon shall fall the slain of all the earth.

bylon ut caderent occisi in Israhel sic de Babylone cadent oc-

51:50 Ye that have escaped the sword, go away, stand cisi in uniuersa terra 51:50 qui fugistis gladium uenite no-

not still: remember the LORD afar off, and let Jerusalem come into your mind.

lite stare recordamini procul Domini et Hierusalem ascendat

51:51 We are confounded, because we have heard re- super cor uestrum 51:51 confusi sumus quoniam audiuimus

proach: shame hath covered our faces: for strangers are come into the sanctuaries of the LORD’s house.

obprobrium operuit ignominia facies nostras quia uenerunt

51:52 Wherefore, behold, the days come, saith the LORD, alieni super sanctificationem domus Domini 51:52 propte-

that I will do judgment upon her graven images: and through all her land the wounded shall groan.

rea ecce dies ueniunt ait Dominus et uisitabo super sculp-

51:53 Though Babylon should mount up to heaven, and tilia eius et in omni terra eius mugiet uulneratus 51:53 si

though she should fortify the height of her strength, yet from me shall spoilers come unto her, saith the LORD.

ascenderit Babylon in caelum et firmauerit in excelso robur

51:54 A sound of a cry cometh from Babylon, and great suum a me uenient uastatores eius ait Dominus 51:54 uox

destruction from the land of the Chaldeans:

Prophetia Ieremiae

clamoris de Babylone et contritio magna de terra Chaldeorum 51:55 quoniam uastauit Dominus Babylonem et perdi- 51:55 Because the LORD hath spoiled Babylon, and de-

stroyed out of her the great voice; when her waves do roar like great waters, a noise of their voice is uttered:

dit ex ea uocem magnam et sonabunt fluctus eorum quasi aquae multae dedit sonitum uox eorum 51:56 quia uenit su- 51:56 Because the spoiler is come upon her, even upon

Babylon, and her mighty men are taken, every one of their bows is broken: for the LORD God of recompences shall

per eam id est super Babylonem praedo et adprehensi sunt surely requite. fortes eius et emarcuit arcus eorum quia fortis ultor Dominus reddens retribuet 51:57 et inebriabo principes eius et 51:57 And I will make drunk her princes, and her wise

men, her captains, and her rulers, and her mighty men: and they shall sleep a perpetual sleep, and not wake, saith

sapientes eius duces eius et magistratus eius et fortes eius the King, whose name is the LORD of hosts. et dormient somnum sempiternum et non expergiscentur ait Rex Dominus exercituum nomen eius 51:58 haec dicit Do- 51:58 Thus saith the LORD of hosts; The broad walls of

Babylon shall be utterly broken, and her high gates shall be burned with fire; and the people shall labour in vain,

minus exercituum murus Babylonis ille latissimus suffos- and the folk in the fire, and they shall be weary. sione suffodietur et portae eius excelsae igni conburentur et labores populorum ad nihilum et gentium in igne erunt et disperibunt 51:59 uerbum quod praecepit Hieremias pro- 51:59 The word which Jeremiah the prophet commanded

Seraiah the son of Neriah, the son of Maaseiah, when he went with Zedekiah the king of Judah into Babylon in the fourth year of his reign. And this Seraiah was a quiet prince.

phetes Saraiae filio Neriae filii Maasiae cum pergeret cum Sedecia rege in Babylonem in anno quarto regni eius Saraias autem erat princeps prophetiae 51:60 et scripsit Hieremias 51:60 So Jeremiah wrote in a book all the evil that should

come upon Babylon, even all these words that are written against Babylon.

omne malum quod uenturum erat super Babylonem in libro uno omnia uerba haec quae scripta sunt contra Babylonem 51:61 et dixit Hieremias ad Saraiam cum ueneris Babylonem 51:61 And Jeremiah said to Seraiah, When thou comest to

Babylon, and shalt see, and shalt read all these words;

et uideris et legeris omnia uerba haec 51:62 dices Domine 51:62 Then shalt thou say, O LORD, thou hast spoken

against this place, to cut it off, that none shall remain in it, neither man nor beast, but that it shall be desolate for

tu locutus es contra locum istum ut disperderes eum ne sit ever. qui in eo habitet ab homine usque ad pecus et ut sit perpetua solitudo 51:63 cumque conpleueris legere librum istum 51:63 And it shall be, when thou hast made an end of read-

ing this book, that thou shalt bind a stone to it, and cast it into the midst of Euphrates:

ligabis ad eum lapidem et proicies illum in medio Eufraten 51:64 et dices sic submergetur Babylon et non consurget a fa- 51:64 And thou shalt say, Thus shall Babylon sink, and

shall not rise from the evil that I will bring upon her: and they shall be weary. Thus far are the words of Jeremiah.

cie adflictionis quam ego adduco super eam et dissoluentur hucusque uerba Hieremiae

52:1 filius uiginti et unius anni Sedecias cum regnare coe- 52:1 Zedekiah was one and twenty years old when he be-

gan to reign, and he reigned eleven years in Jerusalem. And his mother’s name was Hamutal the daughter of

pisset et undecim annis regnauit in Hierusalem et nomen Jeremiah of Libnah. matris eius Amithal filia Hieremiae de Lobna 52:2 et fecit 52:2 And he did that which was evil in the eyes of the

LORD, according to all that Jehoiakim had done.

malum in oculis Domini iuxta omnia quae fecerat Ioachim 52:3 quoniam furor Domini erat in Hierusalem et in Iuda us- 52:3 For through the anger of the LORD it came to pass

in Jerusalem and Judah, till he had cast them out from his presence, that Zedekiah rebelled against the king of Babylon.

Biblia Sacra Vulgata

quequo proiceret eos a facie sua et recessit Sedecias a rege

52:4 And it came to pass in the ninth year of his reign, in Babylonis 52:4 factum est autem in anno nono regni eius the tenth month, in the tenth day of the month, that Nebuchadrezzar king of Babylon came, he and all his army, against Jerusalem, and pitched against it, and built forts against it round about.

in mense decimo decima mensis uenit Nabuchodonosor rex Babylonis ipse et omnis exercitus eius aduersum Hierusalem et obsederunt eam et aedificauerunt contra eam munitiones

52:5 So the city was besieged unto the eleventh year of in circuitu 52:5 et fuit ciuitas obsessa usque ad undecimum

king Zedekiah.

52:6 And in the fourth month, in the ninth day of the annum regis Sedeciae 52:6 mense autem quarto nona men-

month, the famine was sore in the city, so that there was no bread for the people of the land.

sis obtinuit fames in ciuitate et non erant alimenta populo

52:7 Then the city was broken up, and all the men of war terrae 52:7 et disrupta est ciuitas et omnes uiri bellatores

fled, and went forth out of the city by night by the way of the gate between the two walls, which was by the king’s garden; (now the Chaldeans were by the city round about:) and they went by the way of the plain.

fugerunt et exierunt de ciuitate nocte per uiam portae quae est inter duos muros et ducit ad hortum regis Chaldeis obsidentibus urbem in gyro et abierunt per uiam quae ducit

52:8 But the army of the Chaldeans pursued after the king, in heremum 52:8 persecutus est autem exercitus Chaldeoand overtook Zedekiah in the plains of Jericho; and all his army was scattered from him.

rum regem et adprehenderunt Sedeciam in deserto quod est

52:9 Then they took the king, and carried him up unto the iuxta Hiericho et omnis comitatus eius diffugit ab eo 52:9

king of Babylon to Riblah in the land of Hamath; where he gave judgment upon him.

cumque conprehendissent regem adduxerunt eum ad regem Babylonis in Reblatha quae est in terra Emath et locutus est

52:10 And the king of Babylon slew the sons of Zedekiah ad eum iudicia 52:10 et iugulauit rex Babylonis filios Se-

before his eyes: he slew also all the princes of Judah in Riblah.

deciae in oculis eius sed et omnes principes Iudae occidit in

52:11 Then he put out the eyes of Zedekiah; and the king Reblatha 52:11 et oculos Sedeciae eruit et uinxit eum con-

of Babylon bound him in chains, and carried him to Babylon, and put him in prison till the day of his death.

pedibus et adduxit eum rex Babylonis in Babylonem et po-

52:12 Now in the fifth month, in the tenth day of the suit eum in domo carceris usque ad diem mortis eius 52:12

month, which was the nineteenth year of Nebuchadrezzar king of Babylon, came Nebuzaradan, captain of the guard,

which served the king of Babylon, into Jerusalem, in mense autem quinto decima mensis ipse est annus nonusdecimus Nabuchodonosor regis Babylonis uenit Nabuzardan princeps militiae qui stabat coram rege Babylonis in Hieru-

52:13 And burned the house of the LORD, and the king’s salem 52:13 et incendit domum Domini et domum regis et

house; and all the houses of Jerusalem, and all the houses of the great men, burned he with fire:

omnes domos Hierusalem et omnem domum magnam igne

52:14 And all the army of the Chaldeans, that were with conbusit 52:14 et totum murum Hierusalem per circuitum

the captain of the guard, brake down all the walls of Jerusalem round about.

destruxit cunctus exercitus Chaldeorum qui erat cum ma-

52:15 Then Nebuzaradan the captain of the guard carried gistro militiae 52:15 de pauperibus autem populi et de re-

away captive certain of the poor of the people, and the residue of the people that remained in the city, and those that fell away, that fell to the king of Babylon, and the rest of the multitude.

liquo uulgo quod remanserat in ciuitate et de perfugis qui transfugerant ad regem Babylonis et ceteros de multitudine

52:16 But Nebuzaradan the captain of the guard left cer- transtulit Nabuzardan princeps militiae 52:16 de pauperi-

tain of the poor of the land for vinedressers and for husbandmen.

Prophetia Ieremiae

bus uero terrae reliquit Nabuzardan princeps militiae in uinitores et in agricolas 52:17 columnas quoque aereas quae 52:17 Also the pillars of brass that were in the house of

the LORD, and the bases, and the brasen sea that was in the house of the LORD, the Chaldeans brake, and carried

erant in domo Domini et bases et mare aereum quod erat in all the brass of them to Babylon. domo Domini confregerunt Chaldei et tulerunt omne aes eorum in Babylonem 52:18 et lebetas et creagras et psalteria 52:18 The caldrons also, and the shovels, and the snuffers,

and the bowls, and the spoons, and all the vessels of brass wherewith they ministered, took they away.

et fialas et mortariola et omnia uasa aerea quae in ministerio fuerant tulerunt 52:19 et hydrias et thymiamateria et urceos 52:19 And the basons, and the firepans, and the bowls, and

the caldrons, and the candlesticks, and the spoons, and the cups; that which was of gold in gold, and that which was

et pelues et candelabra et mortaria et cyatos quotquot aurea of silver in silver, took the captain of the guard away. aurea et quotquot argentea argentea tulit magister militiae 52:20 columnas duas et mare unum uitulos duodecim aereos 52:20 The two pillars, one sea, and twelve brasen bulls

that were under the bases, which king Solomon had made in the house of the LORD: the brass of all these vessels

qui erant sub basibus quas fecerat rex Salomon in domo Do- was without weight.

mini non erat pondus aeris omnium uasorum horum 52:21 52:21 And concerning the pillars, the height of one pillar was eighteen cubits; and a fillet of twelve cubits did compass it; and the thickness thereof was four fingers: it was

de columnis autem decem et octo cubiti altitudinis erant in hollow. columna una et funiculus duodecim cubitorum circuibat eam porro grossitudo eius quattuor digitorum et intrinsecus caua erat 52:22 et capitella super utramque aerea altitudo capi- 52:22 And a chapiter of brass was upon it; and the

height of one chapiter was five cubits, with network and pomegranates upon the chapiters round about, all of brass. The second pillar also and the pomegranates were like unto these.

telli unius quinque cubitorum et retiacula et mala granata 52:23 nonaginta sex dependentia omnia mala granata cen- 52:23 And there were ninety and six pomegranates on a

side; and all the pomegranates upon the network were an hundred round about.

tum retiaculis circumdabantur 52:24 et tulit magister mili-

52:24 And the captain of the guard took Seraiah the chief priest, and Zephaniah the second priest, and the three keepers of the door:

tiae Saraiam sacerdotem primum et Sophoniam sacerdotem secundum et tres custodes uestibuli 52:25 et de ciuitate tulit 52:25 He took also out of the city an eunuch, which had

the charge of the men of war; and seven men of them that were near the king’s person, which were found in the city; and the principal scribe of the host, who mustered the people of the land; and threescore men of the people of the land, that were found in the midst of the city.

eunuchum unum qui erat praepositus super uiros bellatores et septem uiros de his qui uidebant faciem regis qui inuenti sunt in ciuitate et scribam principem militum qui probabat tirones et sexaginta uiros de populo terrae qui inuenti sunt in medio ciuitatis 52:26 tulit autem eos Nabuzardan magister 52:26 So Nebuzaradan the captain of the guard took them,

and brought them to the king of Babylon to Riblah.

militiae et duxit eos ad regem Babylonis in Reblatha 52:27 52:27 And the king of Babylon smote them, and put them

to death in Riblah in the land of Hamath. Thus Judah was carried away captive out of his own land.

et percussit eos rex Babylonis et interfecit eos in Reblatha in terra Emath et translatus est Iuda de terra sua 52:28 iste 52:28 This is the people whom Nebuchadrezzar carried

away captive: in the seventh year three thousand Jews and three and twenty:

est populus quem transtulit Nabuchodonosor in anno septimo Iudaeos tria milia et uiginti tres 52:29 in anno octauo- 52:29 In the eighteenth year of Nebuchadrezzar he carried

away captive from Jerusalem eight hundred thirty and two persons:

decimo Nabuchodonosor de Hierusalem animas octingentas triginta duas 52:30 in anno uicesimo tertio Nabuchodono- 52:30 In the three and twentieth year of Nebuchadrezzar

Nebuzaradan the captain of the guard carried away captive of the Jews seven hundred forty and five persons: all the persons were four thousand and six hundred.

Biblia Sacra Vulgata

sor transtulit Nabuzardan magister militiae Iudaeorum animas septingentas quadraginta quinque omnes ergo animae

52:31 And it came to pass in the seven and thirtieth year quattuor milia sescentae 52:31 et factum est in tricesimo

of the captivity of Jehoiachin king of Judah, in the twelfth month, in the five and twentieth day of the month, that Evilmerodach king of Babylon in the first year of his reign lifted up the head of Jehoiachin king of Judah, and brought him forth out of prison.

septimo anno transmigrationis Ioachim regis Iudae duodecimo mense uicesima quinta mensis eleuauit Euilmerodach rex Babylonis ipso anno regni sui caput Ioachim regis Iudae

52:32 And spake kindly unto him, and set his throne above et eduxit eum de domo carceris 52:32 et locutus est cum eo the throne of the kings that were with him in Babylon,

bona et posuit thronum eius super thronos regum qui erant

52:33 And changed his prison garments: and he did con- post se in Babylone 52:33 et mutauit uestimenta carceris

tinually eat bread before him all the days of his life.

eius et comedebat panem coram eo semper cunctis diebus

52:34 And for his diet, there was a continual diet given uitae suae 52:34 et cibaria eius cibaria perpetua dabantur

him of the king of Babylon, every day a portion until the day of his death, all the days of his life.

ei a rege Babylonis statuta per singulos dies usque ad diem mortis suae cunctis diebus uitae eius

Lamentationes Ieremiae

1:1 ALEPH quomodo sedit sola ciuitas plena populo facta 1:1 How doth the city sit solitary, that was full of people! how is she become as a widow! she that was great among the nations, and princess among the provinces, how is she

est quasi uidua domina gentium princeps prouinciarum facta become tributary! est sub tributo 1:2 BETH plorans plorauit in nocte et la- 1:2 She weepeth sore in the night, and her tears are on her

cheeks: among all her lovers she hath none to comfort her: all her friends have dealt treacherously with her, they are

crimae eius in maxillis eius non est qui consoletur eam ex become her enemies. omnibus caris eius omnes amici eius spreuerunt eam et facti sunt ei inimici 1:3 GIMEL migrauit Iuda propter adflictio- 1:3 Judah is gone into captivity because of affliction, and

because of great servitude: she dwelleth among the heathen, she findeth no rest: all her persecutors overtook her

nem et multitudinem seruitutis habitauit inter gentes nec in- between the straits. uenit requiem omnes persecutores eius adprehenderunt eam inter angustias 1:4 DELETH uiae Sion lugent eo quod non 1:4 The ways of Zion do mourn, because none come to the

solemn feasts: all her gates are desolate: her priests sigh, her virgins are afflicted, and she is in bitterness.

sint qui ueniant ad sollemnitatem omnes portae eius destructae sacerdotes eius gementes uirgines eius squalidae et ipsa oppressa amaritudine 1:5 HE facti sunt hostes eius in capite 1:5 Her adversaries are the chief, her enemies prosper; for

the LORD hath afflicted her for the multitude of her transgressions: her children are gone into captivity before the

inimici illius locupletati sunt quia Dominus locutus est super enemy. eam propter multitudinem iniquitatum eius paruuli eius ducti sunt captiui ante faciem tribulantis 1:6 VAV et egressus est 1:6 And from the daughter of Zion all her beauty is de-

parted: her princes are become like harts that find no pasture, and they are gone without strength before the pursuer.

a filia Sion omnis decor eius facti sunt principes eius uelut arietes non inuenientes pascuam et abierunt absque fortitudine ante faciem subsequentis 1:7 ZAI recordata est Hieru- 1:7 Jerusalem remembered in the days of her affliction and

of her miseries all her pleasant things that she had in the days of old, when her people fell into the hand of the enemy, and none did help her: the adversaries saw her, and did mock at her sabbaths.

salem dierum adflictionis suae et praeuaricationis omnium desiderabilium suorum quae habuerat a diebus antiquis cum caderet populus eius in manu hostili et non esset auxiliator

uiderunt eam hostes et deriserunt sabbata eius 1:8 HETH 1:8 Jerusalem hath grievously sinned; therefore she is removed: all that honoured her despise her, because they have seen her nakedness: yea, she sigheth, and turneth

peccatum peccauit Hierusalem propterea instabilis facta est backward. omnes qui glorificabant eam spreuerunt illam quia uiderunt ignominiam eius ipsa autem gemens et conuersa retrorsum

1:9 TETH sordes eius in pedibus eius nec recordata est fi- 1:9 Her filthiness is in her skirts; she remembereth not her last end; therefore she came down wonderfully: she had no comforter. O LORD, behold my affliction: for the enemy

nis sui deposita est uehementer non habens consolatorem hath magnified himself. uide Domine adflictionem meam quoniam erectus est inimi-

cus 1:10 IOTH manum suam misit hostis ad omnia desi- 1:10 The adversary hath spread out his hand upon all her pleasant things: for she hath seen that the heathen entered into her sanctuary, whom thou didst command that they

derabilia eius quia uidit gentes ingressas sanctuarium suum should not enter into thy congregation. de quibus praeceperas ne intrarent in ecclesiam tuam 1:11 1:11 All her people sigh, they seek bread; they have given

their pleasant things for meat to relieve the soul: see, O LORD, and consider; for I am become vile.

CAPH omnis populus eius gemens et quaerens panem dederunt pretiosa quaeque pro cibo ad refocilandam animam uide

Biblia Sacra Vulgata

1:12 Is it nothing to you, all ye that pass by? behold, and Domine considera quoniam facta sum uilis 1:12 LAMED o see if there be any sorrow like unto my sorrow, which is done unto me, wherewith the LORD hath afflicted me in

the day of his fierce anger. uos omnes qui transitis per uiam adtendite et uidete si est

dolor sicut dolor meus quoniam uindemiauit me ut locutus

1:13 From above hath he sent fire into my bones, and it est Dominus in die irae furoris sui 1:13 MEM de excelso

prevaileth against them: he hath spread a net for my feet, he hath turned me back: he hath made me desolate and

faint all the day. misit ignem in ossibus meis et erudiuit me expandit rete pe-

dibus meis conuertit me retrorsum posuit me desolatam tota

1:14 The yoke of my transgressions is bound by his hand: die maerore confectam 1:14 NVN uigilauit iugum iniquita-

they are wreathed, and come up upon my neck: he hath made my strength to fall, the LORD hath delivered me

into their hands, from whom I am not able to rise up. tum mearum in manu eius conuolutae sunt et inpositae collo

meo infirmata est uirtus mea dedit me Dominus in manu de

1:15 The LORD hath trodden under foot all my mighty qua non potero surgere 1:15 SAMECH abstulit omnes ma-

men in the midst of me: he hath called an assembly against me to crush my young men: the LORD hath trodden the

virgin, the daughter of Judah, as in a winepress. gnificos meos Dominus de medio mei uocauit aduersum me

tempus ut contereret electos meos torcular calcauit Dominus

1:16 For these things I weep; mine eye, mine eye runneth uirgini filiae Iuda 1:16 AIN idcirco ego plorans et oculus

down with water, because the comforter that should relieve my soul is far from me: my children are desolate, because

the enemy prevailed. meus deducens aquam quia longe factus est a me consolator

conuertens animam meam facti sunt filii mei perditi quo-

1:17 Zion spreadeth forth her hands, and there is none niam inualuit inimicus 1:17 FE expandit Sion manus suas

to comfort her: the LORD hath commanded concerning Jacob, that his adversaries should be round about him:

Jerusalem is as a menstruous woman among them. non est qui consoletur eam mandauit Dominus aduersum Ia-

cob in circuitu eius hostes eius facta est Hierusalem quasi

1:18 The LORD is righteous; for I have rebelled against polluta menstruis inter eos 1:18 SADE iustus est Dominus

his commandment: hear, I pray you, all people, and behold my sorrow: my virgins and my young men are gone into

captivity. quia os eius ad iracundiam prouocaui audite obsecro uni-

uersi populi et uidete dolorem meum uirgines meae et iuue-

1:19 I called for my lovers, but they deceived me: my nes mei abierunt in captiuitatem 1:19 COPH uocaui amicos

priests and mine elders gave up the ghost in the city, while they sought their meat to relieve their souls.

meos et ipsi deceperunt me sacerdotes mei et senes mei in urbe consumpti sunt quia quaesierunt cibum sibi ut refoci-

1:20 Behold, O LORD; for I am in distress: my bowels larent animam suam 1:20 RES uide Domine quoniam tri-

are troubled; mine heart is turned within me; for I have grievously rebelled: abroad the sword bereaveth, at home

there is as death. bulor uenter meus conturbatus est subuersum est cor meum

in memet ipsa quoniam amaritudine plena sum foris inter-

1:21 They have heard that I sigh: there is none to comfort fecit gladius et domi mors similis est 1:21 SEN audierunt

me: all mine enemies have heard of my trouble; they are glad that thou hast done it: thou wilt bring the day that

thou hast called, and they shall be like unto me. quia ingemesco ego et non est qui consoletur me omnes in-

imici mei audierunt malum meum laetati sunt quoniam tu fe-

1:22 Let all their wickedness come before thee; and do cisti adduxisti diem consolationis et fient similes mei 1:22

unto them, as thou hast done unto me for all my transgressions: for my sighs are many, and my heart is faint.

THAV ingrediatur omne malum eorum coram te et deuindemia eos sicut uindemiasti me propter omnes iniquitates meas

Lamentationes Ieremiae

multi enim gemitus mei et cor meum maerens

2:1 ALEPH quomodo obtexit caligine in furore suo Do- 2:1 How hath the LORD covered the daughter of Zion with a cloud in his anger, and cast down from heaven unto the earth the beauty of Israel, and remembered not his foot-

minus filiam Sion proiecit de caelo terram inclitam Israhel stool in the day of his anger! et non recordatus est scabilli pedum suorum in die furoris sui 2:2 BETH praecipitauit Dominus nec pepercit omnia 2:2 The LORD hath swallowed up all the habitations of

Jacob, and hath not pitied: he hath thrown down in his wrath the strong holds of the daughter of Judah; he hath brought them down to the ground: he hath polluted the kingdom and the princes thereof.

speciosa Iacob destruxit in furore suo munitiones uirginis Iuda deiecit in terram polluit regnum et principes eius 2:3 2:3 He hath cut off in his fierce anger all the horn of Israel:

he hath drawn back his right hand from before the enemy, and he burned against Jacob like a flaming fire, which de-

GIMEL confregit in ira furoris omne cornu Israhel auertit re- voureth round about. trorsum dexteram suam a facie inimici et succendit in Iacob

quasi ignem flammae deuorantis in gyro 2:4 DELETH te- 2:4 He hath bent his bow like an enemy: he stood with his right hand as an adversary, and slew all that were pleasant to the eye in the tabernacle of the daughter of Zion: he

tendit arcum suum quasi inimicus firmauit dexteram suam poured out his fury like fire. quasi hostis et occidit omne quod pulchrum erat uisu in tabernaculo filiae Sion effudit quasi ignem indignationem suam 2:5 HE factus est Dominus uelut inimicus praecipita- 2:5 The LORD was as an enemy: he hath swallowed up

Israel, he hath swallowed up all her palaces: he hath destroyed his strong holds, and hath increased in the daugh-

uit Israhel praecipitauit omnia moenia eius dissipauit muni- ter of Judah mourning and lamentation. tiones eius et repleuit in filia Iuda humiliatum et humiliatam

2:6 VAV et dissipauit quasi hortum tentorium suum demo- 2:6 And he hath violently taken away his tabernacle, as if it were of a garden: he hath destroyed his places of the assembly: the LORD hath caused the solemn feasts and sabbaths to be forgotten in Zion, and hath despised in the indignation of his anger the king and the priest.

litus est tabernaculum suum obliuioni tradidit Dominus in Sion festiuitatem et sabbatum et obprobrio in indignatione furoris sui regem et sacerdotem 2:7 ZAI reppulit Dominus 2:7 The LORD hath cast off his altar, he hath abhorred his

sanctuary, he hath given up into the hand of the enemy the walls of her palaces; they have made a noise in the house

altare suum maledixit sanctificationi suae tradidit in manu of the LORD, as in the day of a solemn feast. inimici muros turrium eius uocem dederunt in domo Domini sicut in die sollemni 2:8 HETH cogitauit Dominus dissi- 2:8 The LORD hath purposed to destroy the wall of the

daughter of Zion: he hath stretched out a line, he hath not withdrawn his hand from destroying: therefore he made the rampart and the wall to lament; they languished together.

pare murum filiae Sion tetendit funiculum suum et non auertit manum suam a perditione luxitque antemurale et murus pariter dissipatus est 2:9 TETH defixae sunt in terra portae 2:9 Her gates are sunk into the ground; he hath destroyed

and broken her bars: her king and her princes are among the Gentiles: the law is no more; her prophets also find no

eius perdidit et contriuit uectes eius regem eius et principes vision from the LORD. eius in gentibus non est lex et prophetae eius non inuenerunt uisionem a Domino 2:10 IOTH sederunt in terra con- 2:10 The elders of the daughter of Zion sit upon the

ground, and keep silence: they have cast up dust upon their heads; they have girded themselves with sackcloth: the

ticuerunt senes filiae Sion consperserunt cinere capita sua virgins of Jerusalem hang down their heads to the ground. accincti sunt ciliciis abiecerunt in terra capita sua uirgines Hierusalem 2:11 CAPH defecerunt prae lacrimis oculi mei 2:11 Mine eyes do fail with tears, my bowels are troubled,

my liver is poured upon the earth, for the destruction of the daughter of my people; because the children and the

conturbata sunt uiscera mea effusum est in terra iecur meum sucklings swoon in the streets of the city.

Biblia Sacra Vulgata

super contritione filiae populi mei cum deficeret paruulus et

2:12 They say to their mothers, Where is corn and wine? lactans in plateis oppidi 2:12 LAMED matribus suis dixe-

when they swooned as the wounded in the streets of the city, when their soul was poured out into their mothers’

bosom. runt ubi est triticum et uinum cum deficerent quasi uulne-

rati in plateis ciuitatis cum exhalarent animas suas in sinu

2:13 What thing shall I take to witness for thee? what thing matrum suarum 2:13 MEM cui conparabo te uel cui adsishall I liken to thee, O daughter of Jerusalem? what shall I equal to thee, that I may comfort thee, O virgin daughter of Zion? for thy breach is great like the sea: who can heal thee?

milabo te filia Hierusalem cui exaequabo te et consolabor te uirgo filia Sion magna enim uelut mare contritio tua quis

2:14 Thy prophets have seen vain and foolish things for medebitur tui 2:14 NVN prophetae tui uiderunt tibi falsa et

thee: and they have not discovered thine iniquity, to turn away thy captivity; but have seen for thee false burdens

and causes of banishment. stulta nec aperiebant iniquitatem tuam ut te ad paenitentiam

prouocarent uiderunt autem tibi adsumptiones falsas et eiec-

2:15 All that pass by clap their hands at thee; they hiss and tiones 2:15 SAMECH plauserunt super te manibus omnes wag their head at the daughter of Jerusalem, saying, Is this the city that men call The perfection of beauty, The joy of

the whole earth? transeuntes per uiam sibilauerunt et mouerunt caput suum

super filiam Hierusalem haecine est urbs dicentes perfecti

2:16 All thine enemies have opened their mouth against decoris gaudium uniuersae terrae 2:16 FE aperuerunt su-

thee: they hiss and gnash the teeth: they say, We have swallowed her up: certainly this is the day that we looked

for; we have found, we have seen it. per te os suum omnes inimici tui sibilauerunt et fremuerunt

dentibus dixerunt deuorauimus en ista est dies quam expec-

2:17 The LORD hath done that which he had devised; he tabamus inuenimus uidimus 2:17 AIN fecit Dominus quae

hath fulfilled his word that he had commanded in the days of old: he hath thrown down, and hath not pitied: and he hath caused thine enemy to rejoice over thee, he hath set up the horn of thine adversaries.

cogitauit conpleuit sermonem suum quem praeceperat a diebus antiquis destruxit et non pepercit et laetificauit super te

2:18 Their heart cried unto the LORD, O wall of the inimicum et exaltauit cornu hostium tuorum 2:18 SADE

daughter of Zion, let tears run down like a river day and night: give thyself no rest; let not the apple of thine eye

cease. clamauit cor eorum ad Dominum super muros filiae Sion

deduc quasi torrentem lacrimas per diem et per noctem non

2:19 Arise, cry out in the night: in the beginning of the des requiem tibi neque taceat pupilla oculi tui 2:19 COPH

watches pour out thine heart like water before the face of the LORD: lift up thy hands toward him for the life of thy young children, that faint for hunger in the top of every street.

consurge lauda in nocte in principio uigiliarum effunde sicut aqua cor tuum ante conspectum Domini leua ad eum manus tuas pro anima paruulorum tuorum qui defecerunt in fame

2:20 Behold, O LORD, and consider to whom thou hast in capite omnium conpetorum 2:20 RES uide Domine et

done this. Shall the women eat their fruit, and children of a span long? shall the priest and the prophet be slain in the

sanctuary of the Lord? considera quem uindemiaueris ita ergone comedent mulie-

res fructum suum paruulos ad mensuram palmae si occidetur

2:21 The young and the old lie on the ground in the streets: in sanctuario Domini sacerdos et propheta 2:21 SEN iacue-

my virgins and my young men are fallen by the sword; thou hast slain them in the day of thine anger; thou hast

killed, and not pitied. runt in terra foris puer et senex uirgines meae et iuuenes mei

ceciderunt in gladio interfecisti in die furoris tui percussisti

2:22 Thou hast called as in a solemn day my terrors nec misertus es 2:22 THAV uocasti quasi ad diem sollem-

round about, so that in the day of the LORD’s anger none escaped nor remained: those that I have swaddled and brought up hath mine enemy consumed.

Lamentationes Ieremiae

nem qui terrerent me de circuitu et non fuit in die furoris Domini qui effugeret et relinqueretur quos educaui et enutriui inimicus meus consumpsit eos

3:1 ALEPH ego uir uidens paupertatem meam in uirga in- 3:1 I AM the man that hath seen affliction by the rod of his

wrath.

dignationis eius 3:2 ALEPH me minauit et adduxit in te- 3:2 He hath led me, and brought me into darkness, but not

into light.

nebris et non in lucem 3:3 ALEPH tantum in me uertit 3:3 Surely against me is he turned; he turneth his hand

against me all the day.

et conuertit manum suam tota die 3:4 BETH uetustam fe- 3:4 My flesh and my skin hath he made old; he hath broken

my bones.

cit pellem meam et carnem meam contriuit ossa mea 3:5 3:5 He hath builded against me, and compassed me with

gall and travail.

BETH aedificauit in gyro meo et circumdedit me felle et labore 3:6 BETH in tenebrosis conlocauit me quasi mortuos 3:6 He hath set me in dark places, as they that be dead of

old.

sempiternos 3:7 GIMEL circumaedificauit aduersum me 3:7 He hath hedged me about, that I cannot get out: he

hath made my chain heavy.

ut non egrediar adgrauauit conpedem meam 3:8 GIMEL 3:8 Also when I cry and shout, he shutteth out my prayer.

sed et cum clamauero et rogauero exclusit orationem meam 3:9 GIMEL conclusit uias meas lapidibus quadris semitas 3:9 He hath inclosed my ways with hewn stone, he hath

made my paths crooked.

meas subuertit 3:10 DELETH ursus insidians factus est 3:10 He was unto me as a bear lying in wait, and as a lion

in secret places.

mihi leo in absconditis 3:11 DELETH semitas meas sub- 3:11 He hath turned aside my ways, and pulled me in

pieces: he hath made me desolate.

uertit et confregit me posuit me desolatam 3:12 DELETH 3:12 He hath bent his bow, and set me as a mark for the

arrow.

tetendit arcum suum et posuit me quasi signum ad sagittam 3:13 HE misit in renibus meis filias faretrae suae 3:14 HE 3:13 He hath caused the arrows of his quiver to enter into

my reins.

3:14 I was a derision to all my people; and their song all the day.

factus sum in derisu omni populo meo canticum eorum tota die 3:15 HE repleuit me amaritudinibus inebriauit me ab- 3:15 He hath filled me with bitterness, he hath made me

drunken with wormwood.

sinthio 3:16 VAV et fregit ad numerum dentes meos ciba- 3:16 He hath also broken my teeth with gravel stones, he

hath covered me with ashes.

uit me cinere 3:17 VAV et repulsa est anima mea oblitus 3:17 And thou hast removed my soul far off from peace: I

forgat prosperity.

sum bonorum 3:18 VAV et dixi periit finis meus et spes 3:18 And I said, My strength and my hope is perished from

the LORD:

mea a Domino 3:19 ZAI recordare paupertatis et transgres- 3:19 Remembering mine affliction and my misery, the

wormwood and the gall.

sionis meae absinthii et fellis 3:20 ZAI memoria memor 3:20 My soul hath them still in remembrance, and is hum-

bled in me.

ero et tabescet in me anima mea 3:21 ZAI hoc recolens 3:21 This I recall to my mind, therefore have I hope. in corde meo ideo sperabo 3:22 HETH misericordiae Do- 3:22 It is of the LORD’s mercies that we are not con-

sumed, because his compassions fail not.

mini quia non sumus consumpti quia non defecerunt miserationes eius 3:23 HETH nouae diluculo multa est fides tua 3:23 They are new every morning: great is thy faithful-

ness.

3:24 HETH pars mea Dominus dixit anima mea propterea 3:24 The LORD is my portion, saith my soul; therefore

will I hope in him.

expectabo eum 3:25 TETH bonus est Dominus sperantibus 3:25 The LORD is good unto them that wait for him, to

the soul that seeketh him.

in eum animae quaerenti illum 3:26 TETH bonum est prae- 3:26 It is good that a man should both hope and quietly

wait for the salvation of the LORD.

3:27 It is good for a man that he bear the yoke of his youth. stolari cum silentio salutare Domini 3:27 TETH bonum est

3:28 He sitteth alone and keepeth silence, because he hath uiro cum portauerit iugum ab adulescentia sua 3:28 IOTH borne it upon him.

3:29 He putteth his mouth in the dust; if so be there may sedebit solitarius et tacebit quia leuauit super se 3:29 IOTH

be hope.

3:30 He giveth his cheek to him that smiteth him: he is ponet in puluere os suum si forte sit spes 3:30 IOTH dabit

filled full with reproach.

3:31 For the LORD will not cast off for ever: percutienti se maxillam saturabitur obprobriis 3:31 CAPH

3:32 But though he cause grief, yet will he have compas- quia non repellet in sempiternum Dominus 3:32 CAPH sion according to the multitude of his mercies.

quia si abiecit et miserebitur secundum multitudinem mi-

3:33 For he doth not afflict willingly nor grieve the chil- sericordiarum suarum 3:33 CAPH non enim humiliauit ex dren of men. 3:34 To crush under his feet all the prisoners of the earth. corde suo et abiecit filios hominis 3:34 LAMED ut contere-

3:35 To turn aside the right of a man before the face of the ret sub pedibus suis omnes uinctos terrae 3:35 LAMED ut most High,

3:36 To subvert a man in his cause, the LORD approveth declinaret iudicium uiri in conspectu uultus Altissimi 3:36

not.

LAMED ut peruerteret hominem in iudicio suo Dominus ig-

3:37 Who is he that saith, and it cometh to pass, when the norauit 3:37 MEM quis est iste qui dixit ut fieret Domino Lord commandeth it not?

3:38 Out of the mouth of the most High proceedeth not non iubente 3:38 MEM ex ore Altissimi non egredientur

evil and good? 3:39 Wherefore doth a living man complain, a man for the nec mala nec bona 3:39 MEM quid murmurauit homo uipunishment of his sins?

3:40 Let us search and try our ways, and turn again to the uens uir pro peccatis suis 3:40 NVN scrutemur uias nostras

LORD. 3:41 Let us lift up our heart with our hands unto God in et quaeramus et reuertamur ad Dominum 3:41 NVN leuethe heavens. 3:42 We have transgressed and have rebelled: thou hast mus corda nostra cum manibus ad Dominum in caelos 3:42 not pardoned.

NVN nos inique egimus et ad iracundiam prouocauimus id-

3:43 Thou hast covered with anger, and persecuted us: circo tu inexorabilis es 3:43 SAMECH operuisti in furore et

thou hast slain, thou hast not pitied.

3:44 Thou hast covered thyself with a cloud, that our percussisti nos occidisti nec pepercisti 3:44 SAMECH op-

prayer should not pass through.

3:45 Thou hast made us as the offscouring and refuse in posuisti nubem tibi ne transeat oratio 3:45 SAMECH era-

the midst of the people.

dicationem et abiectionem posuisti me in medio populorum

3:46 All our enemies have opened their mouths against us. 3:46 FE aperuerunt super nos os suum omnes inimici 3:47 3:47 Fear and a snare is come upon us, desolation and destruction.

FE formido et laqueus facta est nobis uaticinatio et contritio

3:48 Mine eye runneth down with rivers of water for the 3:48 FE diuisiones aquarum deduxit oculus meus in contri-

destruction of the daughter of my people.

3:49 Mine eye trickleth down, and ceaseth not, without tione filiae populi mei 3:49 AIN oculus meus adflictus est

any intermission.

3:50 Till the LORD look down, and behold from heaven. nec tacuit eo quod non esset requies 3:50 AIN donec res-3:51 Mine eye affecteth mine heart because of all the piceret et uideret Dominus de caelis 3:51 AIN oculus meus

daughters of my city.

depraedatus est animam meam in cunctis filiabus urbis meae

3:52 Mine enemies chased me sore, like a bird, without 3:52 SADE uenatione ceperunt me quasi auem inimici mei

cause.

3:53 They have cut off my life in the dungeon, and cast a gratis 3:53 SADE lapsa est in lacu uita mea et posuerunt la-

stone upon me.

3:54 Waters flowed over mine head; then I said, I am cut pidem super me 3:54 SADE inundauerunt aquae super ca-

off.

put meum dixi perii 3:55 COPH inuocaui nomen tuum Do- 3:55 I called upon thy name, O LORD, out of the low

dungeon.

mine de lacis nouissimis 3:56 COPH uocem meam audisti 3:56 Thou hast heard my voice: hide not thine ear at my

breathing, at my cry.

ne auertas aurem tuam a singultu meo et clamoribus 3:57 3:57 Thou drewest near in the day that I called upon thee:

thou saidst, Fear not.

COPH adpropinquasti in die quando inuocaui te dixisti ne timeas 3:58 RES iudicasti Domine causam animae meae red- 3:58 O LORD, thou hast pleaded the causes of my soul;

thou hast redeemed my life.

eeptor uitae meae 3:59 RES uidisti Domine iniquitatem ad- 3:59 O LORD, thou hast seen my wrong: judge thou my

cause.

uersum me iudica iudicium meum 3:60 RES uidisti omnem 3:60 Thou hast seen all their vengeance and all their imag-

inations against me.

furorem uniuersas cogitationes eorum aduersum me 3:61 3:61 Thou hast heard their reproach, O LORD, and all

their imaginations against me;

SEN audisti obprobria eorum Domine omnes cogitationes

eorum aduersum me 3:62 SEN labia insurgentium mihi et 3:62 The lips of those that rose up against me, and their

device against me all the day.

meditationes eorum aduersum me tota die 3:63 SEN sessio- 3:63 Behold their sitting down, and their rising up; I am

their musick.

nem eorum et resurrectionem eorum uide ego sum psalmus eorum 3:64 THAV reddes eis uicem Domine iuxta opera 3:64 Render unto them a recompence, O LORD, accord-

ing to the work of their hands.

manuum suarum 3:65 THAV dabis eis scutum cordis labo- 3:65 Give them sorrow of heart, thy curse unto them. rem tuum 3:66 THAV persequeris in furore et conteres eos 3:66 Persecute and destroy them in anger from under the

heavens of the LORD.

sub caelis Domine

4:1 ALEPH quomodo obscuratum est aurum mutatus est 4:1 How is the gold become dim! how is the most fine

gold changed! the stones of the sanctuary are poured out in the top of every street.

color optimus dispersi sunt lapides sanctuarii in capite omnium platearum 4:2 BETH filii Sion incliti et amicti auro 4:2 The precious sons of Zion, comparable to fine gold,

how are they esteemed as earthen pitchers, the work of the hands of the potter!

primo quomodo reputati sunt in uasa testea opus manuum

figuli 4:3 GIMEL sed et lamiae nudauerunt mammam lac- 4:3 Even the sea monsters draw out the breast, they give suck to their young ones: the daughter of my people is become cruel, like the ostriches in the wilderness.

tauerunt catulos suos filia populi mei crudelis quasi strutio

in deserto 4:4 DELETH adhesit lingua lactantis ad pala- 4:4 The tongue of the sucking child cleaveth to the roof of his mouth for thirst: the young children ask bread, and no man breaketh it unto them.

tum eius in siti paruuli petierunt panem et non erat qui fran-

geret eis 4:5 HE qui uescebantur uoluptuose interierunt in 4:5 They that did feed delicately are desolate in the streets:

they that were brought up in scarlet embrace dunghills.

uiis qui nutriebantur in croceis amplexati sunt stercora 4:6 4:6 For the punishment of the iniquity of the daughter of

my people is greater than the punishment of the sin of Sodom, that was overthrown as in a moment, and no hands

VAV et maior effecta est iniquitas filiae populi mei peccato stayed on her. Sodomorum quae subuersa est in momento et non ceperunt in ea manus 4:7 ZAI candidiores nazarei eius niue nitidio- 4:7 Her Nazarites were purer than snow, they were whiter

than milk, they were more ruddy in body than rubies, their polishing was of sapphire:

res lacte rubicundiores ebore antiquo sapphyro pulchriores 4:8 HETH denigrata est super carbones facies eorum et non 4:8 Their visage is blacker than a coal; they are not known

in the streets: their skin cleaveth to their bones; it is withered, it is become like a stick.

sunt cogniti in plateis adhesit cutis eorum ossibus aruit et facta est quasi lignum 4:9 TETH melius fuit occisis gladio 4:9 They that be slain with the sword are better than they

that be slain with hunger: for these pine away, stricken through for want of the fruits of the field.

Biblia Sacra Vulgata

quam interfectis fame quoniam isti extabuerunt consumpti

4:10 The hands of the pitiful women have sodden their ab sterilitate terrae 4:10 IOTH manus mulierum misericor-

own children: they were their meat in the destruction of the daughter of my people.

dium coxerunt filios suos facti sunt cibus earum in contri-

4:11 The LORD hath accomplished his fury; he hath tione filiae populi mei 4:11 CAPH conpleuit Dominus fu-

poured out his fierce anger, and hath kindled a fire in Zion, and it hath devoured the foundations thereof.

rorem suum effudit iram indignationis suae et succendit ig-

4:12 The kings of the earth, and all the inhabitants of the nem in Sion et deuorauit fundamenta eius 4:12 LAMED world, would not have believed that the adversary and the enemy should have entered into the gates of Jerusalem.

non crediderunt reges terrae et uniuersi habitatores orbis quoniam ingrederetur hostis et inimicus per portas Hieru-

4:13 For the sins of her prophets, and the iniquities of her salem 4:13 MEM propter peccata prophetarum eius iniquipriests, that have shed the blood of the just in the midst of her,

tates sacerdotum eius qui effuderunt in medio eius sangui-

4:14 They have wandered as blind men in the streets, they nem iustorum 4:14 NVN errauerunt caeci in plateis polluti

have polluted themselves with blood, so that men could not touch their garments.

sunt sanguine cumque non possent tenuerunt lacinias suas

4:15 They cried unto them, Depart ye; it is unclean; depart, 4:15 SAMECH recedite polluti clamauerunt eis recedite abite

depart, touch not: when they fled away and wandered, they said among the heathen, They shall no more sojourn there.

nolite tangere iurgati quippe sunt et commoti dixerunt in-

4:16 The anger of the LORD hath divided them; he will ter gentes non addet ultra ut habitet in eis 4:16 FE facies

no more regard them: they respected not the persons of the priests, they favoured not the elders.

Domini diuisit eos non addet ut respiciat eos facies sacer-

4:17 As for us, our eyes as yet failed for our vain help: in dotum non erubuerunt neque senum miserti sunt 4:17 AIN our watching we have watched for a nation that could not save us.

cum adhuc subsisteremus defecerunt oculi nostri ad auxilium nostrum uanum cum respiceremus adtenti ad gentem

4:18 They hunt our steps, that we cannot go in our streets: quae saluare non poterat 4:18 SADE lubricauerunt uestigia

our end is near, our days are fulfilled; for our end is come.

nostra in itinere platearum nostrarum adpropinquauit finis

4:19 Our persecutors are swifter than the eagles of the noster conpleti sunt dies nostri quia uenit finis noster 4:19

heaven: they pursued us upon the mountains, they laid wait for us in the wilderness.

COPH uelociores fuerunt persecutores nostri aquilis caeli super montes persecuti sunt nos in deserto insidiati sunt no-

4:20 The breath of our nostrils, the anointed of the LORD, bis 4:20 RES spiritus oris nostri christus dominus captus

was taken in their pits, of whom we said, Under his shadow we shall live among the heathen.

est in peccatis nostris cui diximus in umbra tua uiuemus in

4:21 Rejoice and be glad, O daughter of Edom, that gentibus 4:21 SEN gaude et laetare filia Edom quae habitas

dwellest in the land of Uz; the cup also shall pass through unto thee: thou shalt be drunken, and shalt make thyself

naked. in terra Hus ad te quoque perueniet calix inebriaberis atque

4:22 The punishment of thine iniquity is accomplished, O nudaberis 4:22 THAV conpleta est iniquitas tua filia Sion daughter of Zion; he will no more carry thee away into captivity: he will visit thine iniquity, O daughter of Edom;

he will discover thy sins. non addet ultra ut transmigret te uisitauit iniquitatem tuam filia Edom discoperuit peccata tua

5:1 Remember, O LORD, what is come upon us: consider, 5:1 recordare Domine quid acciderit nobis intuere et re-

and behold our reproach. 5:2 Our inheritance is turned to strangers, our houses to spice obprobrium nostrum 5:2 hereditas nostra uersa est aliens.

ad alienos domus nostrae ad extraneos 5:3 pupilli facti su- 5:3 We are orphans and fatherless, our mothers are as wid-

ows.

mus absque patre matres nostrae quasi uiduae 5:4 aquam 5:4 We have drunken our water for money; our wood is

sold unto us.

nostram pecunia bibimus ligna nostra pretio conparauimus 5:5 ceruicibus minabamur lassis non dabatur requies 5:6 5:5 Our necks are under persecution: we labour, and have

no rest.

5:6 We have given the hand to the Egyptians, and to the Assyrians, to be satisfied with bread.

Aegypto dedimus manum et Assyriis ut saturaremur pane 5:7 patres nostri peccauerunt et non sunt et nos iniquitates 5:7 Our fathers have sinned, and are not; and we have

borne their iniquities.

eorum portauimus 5:8 serui dominati sunt nostri non fuit 5:8 Servants have ruled over us: there is none that doth

deliver us out of their hand.

qui rediieret de manu eorum 5:9 in animabus nostris adfe- 5:9 We gat our bread with the peril of our lives because of

the sword of the wilderness.

rebamus panem nobis a facie gladii in deserto 5:10 pellis 5:10 Our skin was black like an oven because of the terri-

ble famine.

nostra quasi clibanus exusta est a facie tempestatum famis 5:11 mulieres in Sion humiliauerunt uirgines in ciuitatibus 5:11 They ravished the women in Zion, and the maids in

the cities of Judah.

Iuda 5:12 principes manu suspensi sunt facies senum non 5:12 Princes are hanged up by their hand: the faces of

elders were not honoured.

erubuerunt 5:13 adulescentibus inpudice abusi sunt et pueri 5:13 They took the young men to grind, and the children

fell under the wood.

in ligno corruerunt 5:14 senes de portis defecerunt iuuenes 5:14 The elders have ceased from the gate, the young men

from their musick.

de choro psallentium 5:15 defecit gaudium cordis nostri 5:15 The joy of our heart is ceased; our dance is turned

into mourning.

uersus est in luctu chorus noster 5:16 cecidit corona capitis 5:16 The crown is fallen from our head: woe unto us, that

we have sinned!

nostri uae nobis quia peccauimus 5:17 propterea maestum 5:17 For this our heart is faint; for these things our eyes

are dim.

factum est cor nostrum ideo contenebrati sunt oculi nostri 5:18 propter montem Sion quia disperiit uulpes ambulaue- 5:18 Because of the mountain of Zion, which is desolate,

the foxes walk upon it.

runt in eo 5:19 tu autem Domine in aeternum permanebis 5:19 Thou, O LORD, remainest for ever; thy throne from

generation to generation.

solium tuum in generatione et generatione 5:20 quare in 5:20 Wherefore dost thou forget us for ever, and forsake

us so long time?

perpetuum obliuisceris nostri derelinques nos in longitudinem dierum 5:21 conuerte nos Domine ad te et conuerte- 5:21 Turn thou us unto thee, O LORD, and we shall be

turned; renew our days as of old.

mur innoua dies nostros sicut a principio 5:22 sed proiciens 5:22 But thou hast utterly rejected us; thou art very wroth

against us.

reppulisti nos iratus es contra nos uehementer

Biblia Sacra Vulgata

MCLX

Prophetia Ezechielis

1:1 et factum est in tricesimo anno in quarto mense in 1:1 Now it came to pass in the thirtieth year, in the fourth month, in the fifth day of the month, as I was among the captives by the river of Chebar, that the heavens were

quinta mensis cum essem in medio captiuorum iuxta fluuium opened, and I saw visions of God. Chobar aperti sunt caeli et uidi uisiones Dei 1:2 in quinta 1:2 In the fifth day of the month, which was the fifth year

of king Jehoiachin’s captivity,

mensis ipse est annus quintus transmigrationis regis Ioachin 1:3 factum est uerbum Domini ad Hiezecihel filium Buzi sa- 1:3 The word of the LORD came expressly unto Ezekiel

the priest, the son of Buzi, in the land of the Chaldeans by the river Chebar; and the hand of the LORD was there

cerdotem in terra Chaldeorum secus flumen Chobar et facta upon him. est super eum ibi manus Domini 1:4 et uidi et ecce uen- 1:4 And I looked, and, behold, a whirlwind came out of

the north, a great cloud, and a fire infolding itself, and a brightness was about it, and out of the midst thereof as the

tus turbinis ueniebat ab aquilone et nubes magna et ignis colour of amber, out of the midst of the fire. inuoluens et splendor in circuitu eius et de medio eius quasi species electri id est de medio ignis 1:5 et ex medio eorum 1:5 Also out of the midst thereof came the likeness of four

living creatures. And this was their appearance; they had the likeness of a man.

similitudo quattuor animalium et hic aspectus eorum similitudo hominis in eis 1:6 et quattuor facies uni et quattuor 1:6 And every one had four faces, and every one had four

wings.

pinnae uni 1:7 et pedes eorum pedes recti et planta pedis 1:7 And their feet were straight feet; and the sole of their feet was like the sole of a calf’s foot: and they sparkled like the colour of burnished brass.

eorum quasi planta pedis uituli et scintillae quasi aspectus aeris candentis 1:8 et manus hominis sub pinnis eorum in 1:8 And they had the hands of a man under their wings

on their four sides; and they four had their faces and their wings.

quattuor partibus et facies et pinnas per quattuor partes habebant 1:9 iunctaeque erant pinnae eorum alterius ad al- 1:9 Their wings were joined one to another; they turned

not when they went; they went every one straight forward.

terum non reuertebantur cum incederent sed unumquodque ante faciem suam gradiebatur 1:10 similitudo autem uul- 1:10 As for the likeness of their faces, they four had the

face of a man, and the face of a lion, on the right side: and they four had the face of an ox on the left side; they four

tus eorum facies hominis et facies leonis a dextris ipsorum also had the face of an eagle. quattuor facies autem bouis a sinistris ipsorum quattuor et facies aquilae ipsorum quattuor 1:11 et facies eorum et pin- 1:11 Thus were their faces: and their wings were stretched

upward; two wings of every one were joined one to another, and two covered their bodies.

nae eorum extentae desuper duae pinnae singulorum iungebantur et duae tegebant corpora eorum 1:12 et unumquod- 1:12 And they went every one straight forward: whither

the spirit was to go, they went; and they turned not when they went.

que coram facie sua ambulabat ubi erat impetus spiritus illuc gradiebantur nec reuertebantur cum ambularent 1:13 et 1:13 As for the likeness of the living creatures, their ap-

pearance was like burning coals of fire, and like the appearance of lamps: it went up and down among the living creatures; and the fire was bright, and out of the fire went forth lightning.

similitudo animalium aspectus eorum quasi carbonum ignis ardentium et quasi aspectus lampadarum haec erat uisio discurrens in medio animalium splendor ignis et de igne fulgor egrediens 1:14 et animalia ibant et reuertebantur in 1:14 And the living creatures ran and returned as the ap-

pearance of a flash of lightning.

similitudinem fulguris coruscantis 1:15 cumque aspicerem 1:15 Now as I beheld the living creatures, behold one

wheel upon the earth by the living creatures, with his four faces.

animalia apparuit rota una super terram iuxta animalia habens quattuor facies 1:16 et aspectus rotarum et opus ea- 1:16 The appearance of the wheels and their work was like

unto the colour of a beryl: and they four had one likeness: and their appearance and their work was as it were a wheel in the middle of a wheel.

Biblia Sacra Vulgata

rum quasi uisio maris et una similitudo ipsarum quattuor et

1:17 When they went, they went upon their four sides: and aspectus earum et opera quasi sit rota in medio rotae 1:17

they turned not when they went.

per quattuor partes earum euntes ibant et non reuertebantur

1:18 As for their rings, they were so high that they were cum ambularent 1:18 statura quoque erat rotis et altitudo et

dreadful; and their rings were full of eyes round about them four.

horribilis aspectus et totum corpus plenum oculis in circuitu

1:19 And when the living creatures went, the wheels went ipsarum quattuor 1:19 cumque ambularent animalia ambu-

by them: and when the living creatures were lifted up from the earth, the wheels were lifted up.

labant pariter et rotae iuxta ea et cum eleuarentur animalia de

1:20 Whithersoever the spirit was to go, they went, thither terra eleuabantur simul et rotae 1:20 quocumque ibat spiwas their spirit to go; and the wheels were lifted up over against them: for the spirit of the living creature was in the

wheels. ritus illuc eunte spiritu et rotae pariter leuabantur sequentes 1:21 When those went, these went; and when those stood, eum spiritus enim uitae erat in rotis 1:21 cum euntibus ibant

these stood; and when those were lifted up from the earth, the wheels were lifted up over against them: for the spirit

of the living creature was in the wheels. et cum stantibus stabant et cum eleuatis a terra pariter eleua-

bantur et rotae sequentes ea quia spiritus uitae erat in rotis

1:22 And the likeness of the firmament upon the heads of 1:22 et similitudo super caput animalium firmamenti quasi

the living creature was as the colour of the terrible crystal, stretched forth over their heads above.

aspectus cristalli horribilis et extenti super capita eorum de-

1:23 And under the firmament were their wings straight, super 1:23 sub firmamento autem pinnae eorum rectae althe one toward the other: every one had two, which covered on this side, and every one had two, which covered

on that side, their bodies. terius ad alterum unumquodque duabus alis uelabat corpus 1:24 And when they went, I heard the noise of their wings, suum et alterum similiter uelabatur 1:24 et audiebam so-

like the noise of great waters, as the voice of the Almighty, the voice of speech, as the noise of an host: when they

stood, they let down their wings. num alarum quasi sonum aquarum multarum quasi sonum sublimis Dei cum ambularent quasi sonus erat multitudinis ut sonus castrorum cumque starent dimittebantur pinnae eo-

1:25 And there was a voice from the firmament that was rum 1:25 nam cum fieret uox supra firmamentum quod erat

over their heads, when they stood, and had let down their wings.

super caput eorum stabant et submittebant alas suas 1:26 et

1:26 And above the firmament that was over their heads was the likeness of a throne, as the appearance of a sapphire stone: and upon the likeness of the throne was the likeness as the appearance of a man above upon it.

super firmamentum quod erat inminens capiti eorum quasi aspectus lapidis sapphyri similitudo throni et super similitudinem throni similitudo quasi aspectus hominis desuper

1:27 And I saw as the colour of amber, as the appearance 1:27 et uidi quasi speciem electri uelut aspectum ignis intrin-

of fire round about within it, from the appearance of his loins even upward, and from the appearance of his loins even downward, I saw as it were the appearance of fire, and it had brightness round about.

secus eius per circuitum a lumbis eius et desuper et a lumbis eius usque deorsum uidi quasi speciem ignis splendentis in

1:28 As the appearance of the bow that is in the cloud in circuitu 1:28 uelut aspectum arcus cum fuerit in nube in die

the day of rain, so was the appearance of the brightness round about. This was the appearance of the likeness of the glory of the LORD. And when I saw it, I fell upon my face, and I heard a voice of one that spake.

pluuiae hic erat aspectus splendoris per gyrum

2:1 And he said unto me, Son of man, stand upon thy feet, 2:1 haec uisio similitudinis gloriae Domini et uidi et cecidi

and I will speak unto thee.

in faciem meam et audiui uocem loquentis et dixit ad me fili

2:2 And the spirit entered into me when he spake unto me, hominis sta supra pedes tuos et loquar tecum 2:2 et ingres-

and set me upon my feet, that I heard him that spake unto me.

Prophetia Ezechielis

sus est in me spiritus postquam locutus est mihi et statuit me supra pedes meos et audiui loquentem ad me 2:3 et dicen- 2:3 And he said unto me, Son of man, I send thee to

the children of Israel, to a rebellious nation that hath rebelled against me: they and their fathers have transgressed

tem fili hominis mitto ego te ad filios Israhel ad gentes apos- against me, even unto this very day. tatrices quae recesserunt a me patres eorum praeuaricati sunt

pactum meum usque ad diem hanc 2:4 et filii dura facie et 2:4 For they are impudent children and stiffhearted. I do send thee unto them; and thou shalt say unto them, Thus saith the Lord GOD.

indomabili corde sunt ad quos ego mitto te et dices ad eos haec dicit Dominus Deus 2:5 si forte uel ipsi audiant et si 2:5 And they, whether they will hear, or whether they will

forbear, (for they are a rebellious house,) yet shall know that there hath been a prophet among them.

forte quiescant quoniam domus exasperans est et scient quia propheta fuerit in medio eorum 2:6 tu ergo fili hominis ne 2:6 And thou, son of man, be not afraid of them, neither

be afraid of their words, though briers and thorns be with thee, and thou dost dwell among scorpions: be not afraid of their words, nor be dismayed at their looks, though they be a rebellious house.

timeas eos neque sermones eorum metuas quoniam increduli et subuersores sunt tecum et cum scorpionibus habitas uerba eorum ne timeas et uultus eorum ne formides quia domus exasperans est 2:7 loqueris ergo uerba mea ad eos si forte 2:7 And thou shalt speak my words unto them, whether

they will hear, or whether they will forbear: for they are most rebellious.

audiant et quiescant quoniam inritatores sunt 2:8 tu autem

2:8 But thou, son of man, hear what I say unto thee; Be not thou rebellious like that rebellious house: open thy mouth, and eat that I give thee.

fili hominis audi quaecumque loquor ad te et noli esse exasperans sicut domus exasperatrix est aperi os tuum et comede quaecumque ego do tibi 2:9 et uidi et ecce manus missa ad 2:9 And when I looked, behold, an hand was sent unto

me; and, lo, a roll of a book was therein; and he spread it before me; and it was written within and without: and there was written therein lamentations, and mourning, and woe.

me in qua erat inuolutus liber et expandit illum coram me qui erat scriptus intus et foris et scriptae erant in eo lamentationes et carmen et uae

3:1 et dixit ad me fili hominis quodcumque inueneris co- 3:1 Moreover he said unto me, Son of man, eat that thou

findest; eat this roll, and go speak unto the house of Israel.

mede comede uolumen istud et uadens loquere ad filios Israhel 3:2 et aperui os meum et cibauit me uolumine illo 3:3 3:2 So I opened my mouth, and he caused me to eat that

roll.

3:3 And he said unto me, Son of man, cause thy belly to eat, and fill thy bowels with this roll that I give thee. Then did I eat it; and it was in my mouth as honey for sweetness.

et dixit ad me fili hominis uenter tuus comedet et uiscera tua conplebuntur uolumine isto quod ego do tibi et comedi illud et factum est in ore meo sicut mel dulce 3:4 et dixit 3:4 And he said unto me, Son of man, go, get thee unto

the house of Israel, and speak with my words unto them.

ad me fili hominis uade ad domum Israhel et loqueris uerba mea ad eos 3:5 non enim ad populum profundi sermonis et 3:5 For thou art not sent to a people of a strange speech

and of an hard language, but to the house of Israel;

ignotae linguae tu mitteris ad domum Israhel 3:6 neque ad 3:6 Not to many people of a strange speech and of an hard

language, whose words thou canst not understand. Surely, had I sent thee to them, they would have hearkened unto

populos multos profundi sermonis et ignotae linguae quo- thee. rum non possis audire sermones et si ad illos mittereris ipsi audirent te 3:7 domus autem Israhel nolent audire te quia 3:7 But the house of Israel will not hearken unto thee; for

they will not hearken unto me: for all the house of Israel are impudent and hardhearted.

nolunt audire me omnis quippe domus Israhel adtrita fronte

Biblia Sacra Vulgata

3:8 Behold, I have made thy face strong against their faces, est et duro corde 3:8 ecce dedi faciem tuam ualentiorem and thy forehead strong against their foreheads.

faciebus eorum et frontem tuam duriorem frontibus eorum

3:9 As an adamant harder than flint have I made thy fore- 3:9 ut adamantem et ut silicem dedi faciem tuam ne timeas

head: fear them not, neither be dismayed at their looks, though they be a rebellious house.

eos neque metuas a facie eorum quia domus exasperans est

3:10 Moreover he said unto me, Son of man, all my words 3:10 et dixit ad me fili hominis omnes sermones meos quos

that I shall speak unto thee receive in thine heart, and hear with thine ears.

loquor ad te adsume in corde tuo et auribus tuis audi 3:11

3:11 And go, get thee to them of the captivity, unto the children of thy people, and speak unto them, and tell them, Thus saith the Lord GOD; whether they will hear, or whether they will forbear.

et uade ingredere ad transmigrationem ad filios populi tui et loqueris ad eos et dices eis haec dicit Dominus Deus si

3:12 Then the spirit took me up, and I heard behind me a forte audiant et quiescant 3:12 et adsumpsit me spiritus et

voice of a great rushing, saying, Blessed be the glory of the LORD from his place.

audiui post me uocem commotionis magnae benedicta glo-

3:13 I heard also the noise of the wings of the living crea- ria Domini de loco suo 3:13 et uocem alarum animalium tures that touched one another, and the noise of the wheels over against them, and a noise of a great rushing.

percutientium alteram ad alteram et uocem rotarum sequen-

3:14 So the spirit lifted me up, and took me away, and I tium animalia et uocem commotionis magnae 3:14 spiritus

went in bitterness, in the heat of my spirit; but the hand of the LORD was strong upon me.

quoque leuauit me et adsumpsit me et abii amarus in indignatione spiritus mei manus enim Domini erat mecum con-

3:15 Then I came to them of the captivity at Telabib, that fortans me 3:15 et ueni ad transmigrationem aceruum nodwelt by the river of Chebar, and I sat where they sat, and remained there astonished among them seven days.

uarum frugum ad eos qui habitabant iuxta flumen Chobar et sedi ubi illi sedebant et mansi ibi septem diebus maerens

3:16 And it came to pass at the end of seven days, that the in medio eorum 3:16 cum autem pertransissent septem dies

word of the LORD came unto me, saying,

3:17 Son of man, I have made thee a watchman unto the factum est uerbum Domini ad me dicens 3:17 fili hominis

house of Israel: therefore hear the word at my mouth, and give them warning from me.

speculatorem dedi te domui Israhel et audies de ore meo uer-

3:18 When I say unto the wicked, Thou shalt surely die; bum et adnuntiabis eis ex me 3:18 si dicente me ad impium

and thou givest him not warning, nor speakest to warn the wicked from his wicked way, to save his life; the same wicked man shall die in his iniquity; but his blood will I require at thine hand.

morte morieris non adnuntiaueris ei neque locutus fueris ut auertatur a uia sua impia et uiuat ipse impius in iniquitate sua

3:19 Yet if thou warn the wicked, and he turn not from his morietur sanguinem autem eius de manu tua requiram 3:19

wickedness, nor from his wicked way, he shall die in his iniquity; but thou hast delivered thy soul.

si autem tu adnuntiaueris impio et ille non fuerit conuersus ab impietate sua et uia sua impia ipse quidem in iniquitate

3:20 Again, When a righteous man doth turn from his sua morietur tu autem animam tuam liberasti 3:20 sed et si

righteousness, and commit iniquity, and I lay a stumblingblock before him, he shall die: because thou hast not given him warning, he shall die in his sin, and his righteousness which he hath done shall not be remembered; but his blood will I require at thine hand.

conuersus iustus a iustitia sua fecerit iniquitatem ponam offendiculum coram eo ipse morietur quia non adnuntiasti ei in peccato suo morietur et non erunt in memoria iustitiae eius

3:21 Nevertheless if thou warn the righteous man, that the quas fecit sanguinem uero eius de manu tua requiram 3:21 righteous sin not, and he doth not sin, he shall surely live, because he is warned; also thou hast delivered thy soul.

si autem tu adnuntiaueris iusto ut non peccet iustus et ille

Prophetia Ezechielis

non peccauerit uiuens uiuet quia adnuntiasti ei et tu animam tuam liberasti 3:22 et facta est super me manus Domini et 3:22 And the hand of the LORD was there upon me; and

he said unto me, Arise, go forth into the plain, and I will there talk with thee.

dixit ad me surgens egredere in campum et ibi loquar tecum 3:23 et surgens egressus sum in campum et ecce ibi gloria 3:23 Then I arose, and went forth into the plain: and,

behold, the glory of the LORD stood there, as the glory which I saw by the river of Chebar: and I fell on my face.

Domini stabat quasi gloria quam uidi iuxta fluuium Chobar et cecidi in faciem meam 3:24 et ingressus est in me spiri- 3:24 Then the spirit entered into me, and set me upon my

feet, and spake with me, and said unto me, Go, shut thyself within thine house.

tus et statuit me super pedes meos et locutus est mihi et dixit ad me ingredere et includere in medio domus tuae 3:25 et 3:25 But thou, O son of man, behold, they shall put bands

upon thee, and shall bind thee with them, and thou shalt not go out among them:

tu fili hominis ecce data sunt super te uincula et ligabunt te in eis et non egredieris in medio eorum 3:26 et linguam 3:26 And I will make thy tongue cleave to the roof of thy

mouth, that thou shalt be dumb, and shalt not be to them a reprover: for they are a rebellious house.

tuam adherescere faciam palato tuo et eris mutus nec quasi uir obiurgans quia domus exasperans est 3:27 cum autem 3:27 But when I speak with thee, I will open thy mouth,

and thou shalt say unto them, Thus saith the Lord GOD; He that heareth, let him hear; and he that forbeareth, let

locutus fuero tibi aperiam os tuum et dices ad eos haec dicit him forbear: for they are a rebellious house. Dominus Deus qui audit audiat et qui quiescit quiescat quia domus exasperans est

4:1 et tu fili hominis sume tibi laterem et pones eum coram 4:1 Thou also, son of man, take thee a tile, and lay it before

thee, and pourtray upon it the city, even Jerusalem:

te et describes in eo ciuitatem Hierusalem 4:2 et ordinabis 4:2 And lay siege against it, and build a fort against it, and

cast a mount against it; set the camp also against it, and set battering rams against it round about.

aduersus eam obsidionem et aedificabis munitiones et conportabis aggerem et dabis contra eam castra et pones arietes in gyro 4:3 et tu sume tibi sartaginem ferream et pones 4:3 Moreover take thou unto thee an iron pan, and set it

for a wall of iron between thee and the city: and set thy face against it, and it shall be besieged, and thou shalt lay

eam murum ferreum inter te et inter ciuitatem et obfirma- siege against it. This shall be a sign to the house of Israel.

bis faciem tuam ad eam et erit in obsidionem et circumdabis eam signum est domui Israhel 4:4 et tu dormies super latus 4:4 Lie thou also upon thy left side, and lay the iniquity

of the house of Israel upon it: according to the number of the days that thou shalt lie upon it thou shalt bear their

tuum sinistrum et pones iniquitates domus Israhel super eo iniquity. numero dierum quibus dormies super illud et adsumes iniquitatem eorum 4:5 ego autem dedi tibi annos iniquitatis 4:5 For I have laid upon thee the years of their iniquity,

according to the number of the days, three hundred and ninety days: so shalt thou bear the iniquity of the house of

eorum numero dierum trecentos et nonaginta dies et portabis Israel. iniquitatem domus Israhel 4:6 et cum conpleueris haec dor- 4:6 And when thou hast accomplished them, lie again on

thy right side, and thou shalt bear the iniquity of the house of Judah forty days: I have appointed thee each day for a

mies super latus tuum dextrum secundo et adsumes iniqui- year. tatem domus Iuda quadraginta diebus diem pro anno diem

inquam pro anno dedi tibi 4:7 et ad obsidionem Hierusa- 4:7 Therefore thou shalt set thy face toward the siege of Jerusalem, and thine arm shall be uncovered, and thou shalt prophesy against it.

lem conuertes faciem tuam et brachium tuum erit exertum et prophetabis aduersus eam 4:8 ecce circumdedi te uinculis 4:8 And, behold, I will lay bands upon thee, and thou shalt

not turn thee from one side to another, till thou hast ended the days of thy siege.

Biblia Sacra Vulgata

et non te conuertes a latere tuo in latus aliud donec conpleas

4:9 Take thou also unto thee wheat, and barley, and beans, dies obsidionis tuae 4:9 et tu sume tibi frumentum et horand lentiles, and millet, and fitches, and put them in one vessel, and make thee bread thereof, according to the number of the days that thou shalt lie upon thy side, three hundred and ninety days shalt thou eat thereof.

deum et fabam et lentem et milium et uiciam et mittes ea in uas unum et facies tibi panes numero dierum quibus dormies super latus tuum trecentis et nonaginta diebus come-

4:10 And thy meat which thou shalt eat shall be by weight, des illud 4:10 cibus autem tuus quo uesceris erit in pondere

twenty shekels a day: from time to time shalt thou eat it.

uiginti stateres in die a tempore usque ad tempus comedes

4:11 Thou shalt drink also water by measure, the sixth part illud 4:11 et aquam in mensura bibes sextam partem hin a of an hin: from time to time shalt thou drink.

4:12 And thou shalt eat it as barley cakes, and thou shalt tempore usque ad tempus bibes illud 4:12 et quasi subcine-

bake it with dung that cometh out of man, in their sight.

ricium hordiacium comedes illud et stercore quod egredietur

4:13 And the LORD said, Even thus shall the children of de homine operies illud in oculis eorum 4:13 et dixit Do-

Israel eat their defiled bread among the Gentiles, whither I will drive them.

minus sic comedent filii Israhel panem suum pollutum inter

4:14 Then said I, Ah Lord GOD! behold, my soul hath not gentes ad quas eiciam eos 4:14 et dixi ha ha ha Domine been polluted: for from my youth up even till now have I not eaten of that which dieth of itself, or is torn in pieces;

neither came there abominable flesh into my mouth. Deus ecce anima mea non est polluta et morticinum et la-

ceratum a bestiis non comedi ab infantia mea usque nunc

4:15 Then he said unto me, Lo, I have given thee cow’s et non est ingressa os meum omnis caro inmunda 4:15 et

dung for man’s dung, and thou shalt prepare thy bread therewith.

dixit ad me ecce dedi tibi fimum boum pro stercoribus hu-

4:16 Moreover he said unto me, Son of man, behold, I manis et facies panem tuum in eo 4:16 et dixit ad me fili

will break the staff of bread in Jerusalem: and they shall eat bread by weight, and with care; and they shall drink

water by measure, and with astonishment: hominis ecce ego conteram baculum panis in Hierusalem et

comedent panem in pondere et in sollicitudine et aquam in

4:17 That they may want bread and water, and be astonied mensura et in angustia bibent 4:17 ut deficientibus pane et

one with another, and consume away for their iniquity.

aqua corruat unusquisque ad fratrem suum et contabescant in iniquitatibus suis

5:1 And thou, son of man, take thee a sharp knife, take 5:1 et tu fili hominis sume tibi gladium acutum radentem

thee a barber’s razor, and cause it to pass upon thine head and upon thy beard: then take thee balances to weigh, and

divide the hair. pilos adsumes eum et duces per caput tuum et per barbam 5:2 Thou shalt burn with fire a third part in the midst of tuam et adsumes tibi stateram ponderis et diuides eos 5:2

the city, when the days of the siege are fulfilled: and thou shalt take a third part, and smite about it with a knife: and a third part thou shalt scatter in the wind; and I will draw out a sword after them.

tertiam partem igni conbures in medio ciuitatis iuxta conpletionem dierum obsidionis et adsumens tertiam partem concides gladio in circuitu eius tertiam uero aliam disperges in

5:3 Thou shalt also take thereof a few in number, and bind uentum et gladium nudabo post eos 5:3 et sumes inde parthem in thy skirts.

5:4 Then take of them again, and cast them into the midst uum numerum et ligabis eos in summitate pallii tui 5:4 et

of the fire, and burn them in the fire; for thereof shall a fire come forth into all the house of Israel.

ex eis rursum tolles et proicies in medio ignis et conbures eos

5:5 Thus saith the Lord GOD; This is Jerusalem: I have set igni ex eo egredietur ignis in omnem domum Israhel 5:5 it in the midst of the nations and countries that are round about her.

Prophetia Ezechielis

haec dicit Dominus Deus ista est Hierusalem in medio gentium posui eam et in circuitu eius terras 5:6 et contempsit 5:6 And she hath changed my judgments into wickedness

more than the nations, and my statutes more than the countries that are round about her: for they have refused my

iudicia mea ut plus esset impia quam gentes et praecepta mea judgments and my statutes, they have not walked in them. ultra quam terrae quae in circuitu eius sunt iudicia enim mea proiecerunt et in praeceptis meis non ambulauerunt 5:7 id- 5:7 Therefore thus saith the Lord GOD; Because ye multi-

plied more than the nations that are round about you, and have not walked in my statutes, neither have kept my judgments, neither have done according to the judgments of the nations that are round about you;

circo haec dicit Dominus Deus quia superastis gentes quae in circuitu uestro sunt in praeceptis meis non ambulastis et iudicia mea non fecistis et iuxta iudicia gentium quae in circuitu uestro sunt non estis operati 5:8 ideo haec dicit Do- 5:8 Therefore thus saith the Lord GOD; Behold, I, even I,

am against thee, and will execute judgments in the midst of thee in the sight of the nations.

minus Deus ecce ego ad te et ipse ego faciam in medio tui iudicia in oculis gentium 5:9 et faciam in te quae non feci 5:9 And I will do in thee that which I have not done, and

whereunto I will not do any more the like, because of all thine abominations.

et quibus similia ultra non faciam propter omnes abomina-

tiones tuas 5:10 ideo patres comedent filios in medio tui 5:10 Therefore the fathers shall eat the sons in the midst of thee, and the sons shall eat their fathers; and I will execute judgments in thee, and the whole remnant of thee will I

et filii comedent patres suos et faciam in te iudicia et uen- scatter into all the winds. tilabo uniuersas reliquias tuas in omnem uentum 5:11 id- 5:11 Wherefore, as I live, saith the Lord GOD; Surely,

because thou hast defiled my sanctuary with all thy detestable things, and with all thine abominations, therefore will I also diminish thee; neither shall mine eye spare, neither will I have any pity.

circo uiuo ego dicit Dominus Deus nisi pro eo quod sanctum meum uiolasti in omnibus offensionibus tuis et in omnibus abominationibus tuis ego quoque confringam et non parcet oculus meus et non miserebor 5:12 tertia tui pars peste mo- 5:12 A third part of thee shall die with the pestilence, and

with famine shall they be consumed in the midst of thee: and a third part shall fall by the sword round about thee; and I will scatter a third part into all the winds, and I will draw out a sword after them.

rietur et fame consumetur in medio tui et tertia tui pars gladio cadet in circuitu tuo tertiam uero partem tuam in omnem uentum dispergam et gladium euaginabo post eos 5:13 et 5:13 Thus shall mine anger be accomplished, and I will

cause my fury to rest upon them, and I will be comforted: and they shall know that I the LORD have spoken it in my

conpleam furorem meum et requiescere faciam indignatio- zeal, when I have accomplished my fury in them. nem meam in eis et consolabor et scient quia ego Dominus locutus sum in zelo meo cum impleuero indignationem meam in eis 5:14 et dabo te in desertum et in obprobrium in 5:14 Moreover I will make thee waste, and a reproach

among the nations that are round about thee, in the sight of all that pass by.

gentibus quae in circuitu tuo sunt in conspectu omnis praetereuntis 5:15 et eris obprobrium et blasphemia exemplum 5:15 So it shall be a reproach and a taunt, an instruction

and an astonishment unto the nations that are round about thee, when I shall execute judgments in thee in anger and

et stupor in gentibus quae in circuitu tuo sunt cum fecero in in fury and in furious rebukes. I the LORD have spoken it. te iudicia in furore et in indignatione et in increpationibus irae 5:16 ego Dominus locutus sum quando misero sagittas 5:16 When I shall send upon them the evil arrows of

famine, which shall be for their destruction, and which I will send to destroy you: and I will increase the famine

famis pessimas in eos quae erunt mortiferae et quas mittam upon you, and will break your staff of bread: ut disperdam uos et famem congregabo super uos et conte-

Biblia Sacra Vulgata

5:17 So will I send upon you famine and evil beasts, and ram uobis baculum panis 5:17 et inmittam in uos famem they shall bereave thee: and pestilence and blood shall pass through thee; and I will bring the sword upon thee.

I the LORD have spoken it. et bestias pessimas usque ad internicionem et pestilentia et sanguis transibunt per te et gladium inducam super te ego Dominus locutus sum

6:1 And the word of the LORD came unto me, saying, 6:1 et factus est sermo Domini ad me dicens 6:2 fili ho-

6:2 Son of man, set thy face toward the mountains of Israel, and prophesy against them,

minis pone faciem tuam ad montes Israhel et prophetabis

6:3 And say, Ye mountains of Israel, hear the word of the ad eos 6:3 et dices montes Israhel audite uerbum Domini Lord GOD; Thus saith the Lord GOD to the mountains, and to the hills, to the rivers, and to the valleys; Behold, I, even I, will bring a sword upon you, and I will destroy your high places.

Dei haec dicit Dominus Deus montibus et collibus rupibus et uallibus ecce ego inducam super uos gladium et disperdam

6:4 And your altars shall be desolate, and your images excelsa uestra 6:4 et demoliar aras uestras et confringen-

shall be broken: and I will cast down your slain men before your idols.

tur simulacra uestra et deiciam interfectos uestros ante idola

6:5 And I will lay the dead carcases of the children of Is- uestra 6:5 et dabo cadauera filiorum Israhel ante faciem si-

rael before their idols; and I will scatter your bones round about your altars.

mulacrorum uestrorum et dispergam ossa uestra circum aras

6:6 In all your dwellingplaces the cities shall be laid waste, uestras 6:6 in omnibus habitationibus uestris urbes deserand the high places shall be desolate; that your altars may be laid waste and made desolate, and your idols may be broken and cease, and your images may be cut down, and your works may be abolished.

tae erunt et excelsa demolientur et dissipabuntur et interibunt arae uestrae et confringentur et cessabunt idola uestra et

6:7 And the slain shall fall in the midst of you, and ye shall conterentur delubra uestra et delebuntur opera uestra 6:7 et

know that I am the LORD.

cadet interfectus in medio uestri et scietis quoniam ego Do-

6:8 Yet will I leave a remnant, that ye may have some that minus 6:8 et relinquam in uobis eos qui fugerint gladium in

shall escape the sword among the nations, when ye shall be scattered through the countries.

gentibus cum dispersero uos in terris 6:9 et recordabuntur

6:9 And they that escape of you shall remember me among the nations whither they shall be carried captives, because I am broken with their whorish heart, which hath departed from me, and with their eyes, which go a whoring after their idols: and they shall lothe themselves for the evils which they have committed in all their abominations.

mei liberati uestri in gentibus ad quas captiui ducti sunt quia contriui cor eorum fornicans et recedens a me et oculos eorum fornicantes post idola sua et displicebunt sibimet super

6:10 And they shall know that I am the LORD, and that I malis quae fecerunt in uniuersis abominationibus suis 6:10

have not said in vain that I would do this evil unto them.

et scient quia ego Dominus non frustra locutus sum ut face-

6:11 Thus saith the Lord GOD; Smite with thine hand, and rem eis malum hoc 6:11 haec dicit Dominus Deus percute stamp with thy foot, and say, Alas for all the evil abominations of the house of Israel! for they shall fall by the

sword, by the famine, and by the pestilence. manu tua et adlide pedem tuum et dic eheu ad omnes abomi-

nationes malorum domus Israhel qui gladio fame peste rui-

6:12 He that is far off shall die of the pestilence; and he turi sunt 6:12 qui longe est peste morietur qui autem prope

that is near shall fall by the sword; and he that remaineth and is besieged shall die by the famine: thus will I accom-

plish my fury upon them. gladio corruet et qui relictus fuerit et obsessus fame morie-6:13 Then shall ye know that I am the LORD, when their tur et conpleam indignationem meam in eis 6:13 et scietis

slain men shall be among their idols round about their altars, upon every high hill, in all the tops of the mountains, and under every green tree, and under every thick oak, the place where they did offer sweet savour to all their idols.

quia ego Dominus cum fuerint interfecti uestri in medio idolorum uestrorum in circuitu ararum uestrarum in omni colle

Prophetia Ezechielis

excelso in cunctis summitatibus montium et subtus omne lignum nemorosum et subtus uniuersam quercum frondosam locum ubi accenderunt tura redolentia uniuersis idolis suis 6:14 et extendam manum meam super eos et faciam terram 6:14 So will I stretch out my hand upon them, and make

the land desolate, yea, more desolate than the wilderness toward Diblath, in all their habitations: and they shall

desolatam et destitutam a deserto Deblatha in omnibus habi- know that I am the LORD. tationibus eorum et scient quia ego Dominus

7:1 et factus est sermo Domini ad me dicens 7:2 et tu 7:1 Moreover the word of the LORD came unto me, say-

ing,

7:2 Also, thou son of man, thus saith the Lord GOD unto the land of Israel; An end, the end is come upon the four corners of the land.

fili hominis haec dicit Dominus Deus terrae Israhel finis uenit finis super quattuor plagas terrae 7:3 nunc finis super 7:3 Now is the end come upon thee, and I will send mine

anger upon thee, and will judge thee according to thy ways, and will recompense upon thee all thine abomina-

te et emittam furorem meum in te et iudicabo te iuxta uias tions.

tuas et ponam contra te omnes abominationes tuas 7:4 et 7:4 And mine eye shall not spare thee, neither will I have pity: but I will recompense thy ways upon thee, and thine abominations shall be in the midst of thee: and ye shall

non parcet oculus meus super te et non miserebor sed uias know that I am the LORD. tuas ponam super te et abominationes tuae in medio tui erunt et scietis quia ego Dominus 7:5 haec dicit Dominus Deus 7:5 Thus saith the Lord GOD; An evil, an only evil, be-

hold, is come.

adflictio una adflictio ecce uenit 7:6 finis uenit uenit finis 7:6 An end is come, the end is come: it watcheth for thee;

behold, it is come.

euigilauit aduersum te ecce uenit 7:7 uenit contractio super 7:7 The morning is come unto thee, O thou that dwellest

in the land: the time is come, the day of trouble is near, and not the sounding again of the mountains.

te qui habitas in terra uenit tempus prope est dies occisionis et non gloriae montium 7:8 nunc de propinquo effun- 7:8 Now will I shortly pour out my fury upon thee, and

accomplish mine anger upon thee: and I will judge thee according to thy ways, and will recompense thee for all

dam iram meam super te et conpleam furorem meum in te thine abominations. et iudicabo te iuxta uias tuas et inponam tibi omnia scelera tua 7:9 et non parcet oculus meus neque miserebor sed uias 7:9 And mine eye shall not spare, neither will I have pity:

I will recompense thee according to thy ways and thine abominations that are in the midst of thee; and ye shall

tuas inponam tibi et abominationes tuae in medio tui erunt know that I am the LORD that smiteth. et scietis quia ego sum Dominus percutiens 7:10 ecce dies 7:10 Behold the day, behold, it is come: the morning is

gone forth; the rod hath blossomed, pride hath budded.

ecce uenit egressa est contractio floruit uirga germinauit superbia 7:11 iniquitas surrexit in uirga impietatis non ex eis 7:11 Violence is risen up into a rod of wickedness: none

of them shall remain, nor of their multitude, nor of any of their’s: neither shall there be wailing for them.

et non ex populo neque ex sonitu eorum et non erit requies in eis 7:12 uenit tempus adpropinquauit dies qui emit non 7:12 The time is come, the day draweth near: let not the

buyer rejoice, nor the seller mourn: for wrath is upon all the multitude thereof.

laetetur et qui uendit non lugeat quia ira super omnem po-

pulum eius 7:13 quia qui uendit ad id quod uendidit non 7:13 For the seller shall not return to that which is sold, although they were yet alive: for the vision is touching the whole multitude thereof, which shall not return; neither

reuertetur et adhuc in uiuentibus uita eorum uisio enim ad shall any strengthen himself in the iniquity of his life. omnem multitudinem eius non regredietur et uir in iniquitate uitae suae non confortabitur 7:14 canite tuba praepa- 7:14 They have blown the trumpet, even to make all ready;

but none goeth to the battle: for my wrath is upon all the multitude thereof.

rentur omnes et non est qui uadat ad proelium ira enim mea

Biblia Sacra Vulgata

7:15 The sword is without, and the pestilence and the super uniuersum populum eius 7:15 gladius foris pestis et

famine within: he that is in the field shall die with the sword; and he that is in the city, famine and pestilence

shall devour him. fames intrinsecus qui in agro est gladio morietur et qui in ci-7:16 But they that escape of them shall escape, and shall uitate pestilentia et fame deuorabuntur 7:16 et saluabuntur

be on the mountains like doves of the valleys, all of them mourning, every one for his iniquity.

qui fugerint ex eis et erunt in montibus quasi columbae con-

7:17 All hands shall be feeble, and all knees shall be weak uallium omnes trepidi unusquisque in iniquitate sua 7:17

as water. 7:18 They shall also gird themselves with sackcloth, and omnes manus dissoluentur et omnia genua fluent aquis 7:18 horror shall cover them; and shame shall be upon all faces, and baldness upon all their heads.

et accingent se ciliciis et operiet eos formido et in omni fa-

7:19 They shall cast their silver in the streets, and their cie confusio et in uniuersis capitibus eorum caluitium 7:19

gold shall be removed: their silver and their gold shall not be able to deliver them in the day of the wrath of the LORD: they shall not satisfy their souls, neither fill their bowels: because it is the stumblingblock of their iniquity.

argentum eorum foris proicietur et aurum eorum in sterquilinium erit argentum eorum et aurum eorum non ualebit liberare eos in die furoris Domini animam suam non saturabunt et uentres eorum non implebuntur quia scandalum iniquita-

7:20 As for the beauty of his ornament, he set it in majesty: tis eorum factum est 7:20 et ornamentum monilium suorum but they made the images of their abominations and of their detestable things therein: therefore have I set it far

from them. in superbiam posuerunt et imagines abominationum suarum

et simulacrorum fecerunt ex eo propter hoc dedi eis illud

7:21 And I will give it into the hands of the strangers for in inmunditiam 7:21 et dabo illud in manus alienorum ad a prey, and to the wicked of the earth for a spoil; and they shall pollute it.

diripiendum et impiis terrae in praedam et contaminabunt

7:22 My face will I turn also from them, and they shall illud 7:22 et auertam faciem meam ab eis et uiolabunt arca-

pollute my secret place: for the robbers shall enter into it, and defile it.

num meum et introibunt in illud emissarii et contaminabunt

7:23 Make a chain: for the land is full of bloody crimes, illud 7:23 fac conclusionem quoniam terra plena est iudi-

and the city is full of violence.

7:24 Wherefore I will bring the worst of the heathen, and cio sanguinum et ciuitas plena iniquitate 7:24 et adducam

they shall possess their houses: I will also make the pomp of the strong to cease; and their holy places shall be de-

filed. pessimos de gentibus et possidebunt domos eorum et quies-

cere faciam superbiam potentium et possidebunt sanctuaria

7:25 Destruction cometh; and they shall seek peace, and eorum 7:25 angustia superueniente requirent pacem et non

there shall be none.

7:26 Mischief shall come upon mischief, and rumour shall erit 7:26 conturbatio super conturbationem ueniet et audi-

be upon rumour; then shall they seek a vision of the prophet; but the law shall perish from the priest, and coun-

sel from the ancients. tus super auditum et quaerent uisionem de propheta et lex 7:27 The king shall mourn, and the prince shall be clothed peribit a sacerdote et consilium a senioribus 7:27 rex luge-

with desolation, and the hands of the people of the land shall be troubled: I will do unto them after their way, and according to their deserts will I judge them; and they shall know that I am the LORD.

bit et princeps induetur maerore et manus populi terrae conturbabuntur secundum uiam eorum faciam eis et secundum iudicia eorum iudicabo eos et scient quia ego Dominus

8:1 And it came to pass in the sixth year, in the sixth 8:1 et factum est in anno sexto in sexto mense in quinta

month, in the fifth day of the month, as I sat in mine house, and the elders of Judah sat before me, that the hand of the

Lord GOD fell there upon me. mensis ego sedebam in domo mea et senes Iuda sedebant 8:2 Then I beheld, and lo a likeness as the appearance coram me et cecidit super me ibi manus Domini Dei 8:2 et

of fire: from the appearance of his loins even downward, fire; and from his loins even upward, as the appearance of brightness, as the colour of amber.

Prophetia Ezechielis

uidi et ecce similitudo quasi aspectus ignis ab aspectu lumborum eius et deorsum ignis et a lumbis eius et sursum quasi aspectus splendoris ut uisio electri 8:3 et emissa similitudo 8:3 And he put forth the form of an hand, and took me by

a lock of mine head; and the spirit lifted me up between the earth and the heaven, and brought me in the visions of God to Jerusalem, to the door of the inner gate that looketh toward the north; where was the seat of the image of jealousy, which provoketh to jealousy.

manus adprehendit me in cincinno capitis mei et eleuauit me spiritus inter terram et caelum et adduxit in Hierusalem in uisione Dei iuxta ostium interius quod respiciebat aquilonem ubi erat statutum idolum zeli ad prouocandam aemulationem 8:4 et ecce ibi gloria Dei Israhel secundum uisionem quam 8:4 And, behold, the glory of the God of Israel was there,

according to the vision that I saw in the plain.

uideram in campo 8:5 et dixit ad me fili hominis leua ocu- 8:5 Then said he unto me, Son of man, lift up thine eyes

now the way toward the north. So I lifted up mine eyes the way toward the north, and behold northward at the gate of

los tuos ad uiam aquilonis et leuaui oculos meos ad uiam the altar this image of jealousy in the entry. aquilonis et ecce ab aquilone portae altaris idolum zeli in ipso introitu 8:6 et dixit ad me fili hominis putasne uides tu 8:6 He said furthermore unto me, Son of man, seest thou

what they do? even the great abominations that the house of Israel committeth here, that I should go far off from my sanctuary? but turn thee yet again, and thou shalt see greater abominations.

quid isti faciant abominationes magnas quas domus Israhel facit hic ut procul recedam a sanctuario meo et adhuc conuersus uidebis abominationes maiores 8:7 et introduxit me 8:7 And he brought me to the door of the court; and when

I looked, behold a hole in the wall.

ad ostium atrii et uidi et ecce foramen unum in pariete 8:8 8:8 Then said he unto me, Son of man, dig now in the wall:

and when I had digged in the wall, behold a door.

et dixit ad me fili hominis fode parietem et cum perfodissem parietem apparuit ostium unum 8:9 et dixit ad me ingredere 8:9 And he said unto me, Go in, and behold the wicked

abominations that they do here.

et uide abominationes pessimas quas isti faciunt hic 8:10 8:10 So I went in and saw; and behold every form of creep-

ing things, and abominable beasts, and all the idols of the house of Israel, pourtrayed upon the wall round about.

et ingressus uidi et ecce omnis similitudo reptilium et animalium abominatio et uniuersa idola domus Israhel depicta erant in pariete in circuitu per totum 8:11 et septuaginta uiri 8:11 And there stood before them seventy men of the an-

cients of the house of Israel, and in the midst of them stood Jaazaniah the son of Shaphan, with every man his censer

de senioribus domus Israhel et Hiezonias filius Saphan stabat in his hand; and a thick cloud of incense went up. in medio eorum stantium ante picturas et unusquisque habebat turibulum in manu sua et uapor nebulae de ture consurgebat 8:12 et dixit ad me certe uides fili hominis quae seniores 8:12 Then said he unto me, Son of man, hast thou seen

what the ancients of the house of Israel do in the dark, every man in the chambers of his imagery? for they say,

domus Israhel faciunt in tenebris unusquisque in abscondito the LORD seeth us not; the LORD hath forsaken the earth. cubiculi sui dicunt enim non uidet Dominus nos dereliquit Dominus terram 8:13 et dixit ad me adhuc conuersus uide- 8:13 He said also unto me, Turn thee yet again, and thou

shalt see greater abominations that they do.

bis abominationes maiores quas isti faciunt 8:14 et intro- 8:14 Then he brought me to the door of the gate of the

LORD’s house which was toward the north; and, behold, there sat women weeping for Tammuz.

duxit me per ostium portae domus Domini quod respiciebat ad aquilonem et ecce ibi mulieres sedebant plangentes Adonidem 8:15 et dixit ad me certe uidisti fili hominis adhuc 8:15 Then said he unto me, Hast thou seen this, O son

of man? turn thee yet again, and thou shalt see greater abominations than these.

Biblia Sacra Vulgata

8:16 And he brought me into the inner court of the conuersus uidebis abominationes maiores his 8:16 et intro-

LORD’s house, and, behold, at the door of the temple of the LORD, between the porch and the altar, were about five and twenty men, with their backs toward the temple of the LORD, and their faces toward the east; and they worshipped the sun toward the east.

duxit me in atrium domus Domini interius et ecce in ostio templi Domini inter uestibulum et altare quasi uiginti quinque uiri dorsa habentes contra templum Domini et facies ad

8:17 Then he said unto me, Hast thou seen this, O son of orientem et adorabant ad ortum solis 8:17 et dixit ad me

man? Is it a light thing to the house of Judah that they commit the abominations which they commit here? for they have filled the land with violence, and have returned to provoke me to anger: and, lo, they put the branch to their nose.

certe uidisti fili hominis numquid leue est hoc domui Iuda ut facerent abominationes istas quas fecerunt hic quia replentes terram iniquitate conuersi sunt ad inritandum me et ecce

8:18 Therefore will I also deal in fury: mine eye shall not adplicant ramum ad nares suas 8:18 ergo et ego faciam in spare, neither will I have pity: and though they cry in mine ears with a loud voice, yet will I not hear them.

furore non parcet oculus meus nec miserebor et cum clamauerint ad aures meas uoce magna non exaudiam eos

9:1 He cried also in mine ears with a loud voice, saying, 9:1 et clamauit in auribus meis uoce magna dicens adpro-

Cause them that have charge over the city to draw near, even every man with his destroying weapon in his hand.

pinquauerunt uisitationes urbis et unusquisque uas interfec-

9:2 And, behold, six men came from the way of the higher tionis habet in manu sua 9:2 et ecce sex uiri ueniebant de

gate, which lieth toward the north, and every man a slaughter weapon in his hand; and one man among them was clothed with linen, with a writer’s inkhorn by his side: and they went in, and stood beside the brasen altar.

uia portae superioris quae respicit ad aquilonem et uniuscuiusque uas interitus in manu eius uir quoque unus in medio eorum uestitus lineis et atramentarium scriptoris ad renes

9:3 And the glory of the God of Israel was gone up from eius et ingressi sunt et steterunt iuxta altare aereum 9:3 et

the cherub, whereupon he was, to the threshold of the house. And he called to the man clothed with linen, which

had the writer’s inkhorn by his side; gloria Domini Israhel adsumpta est de cherub quae erat su-

per eum ad limen domus et uocauit uirum qui indutus erat li-

9:4 And the LORD said unto him, Go through the midst neis et atramentarium scriptoris habebat in lumbis suis 9:4

of the city, through the midst of Jerusalem, and set a mark upon the foreheads of the men that sigh and that cry for all

the abominations that be done in the midst thereof. et dixit Dominus ad eum transi per mediam ciuitatem in medio Hierusalem et signa thau super frontes uirorum gementium et dolentium super cunctis abominationibus quae fiunt

9:5 And to the others he said in mine hearing, Go ye after in medio eius 9:5 et illis dixit audiente me transite per ci-

him through the city, and smite: let not your eye spare, neither have ye pity:

uitatem sequentes eum et percutite non parcat oculus ues-

9:6 Slay utterly old and young, both maids, and little chil- ter neque misereamini 9:6 senem adulescentulum et uirgidren, and women: but come not near any man upon whom is the mark; and begin at my sanctuary. Then they began

at the ancient men which were before the house. nem paruulum et mulieres interficite usque ad internicionem omnem autem super quem uideritis thau ne occidatis et a sanctuario meo incipite coeperunt ergo a uiris senioribus qui

9:7 And he said unto them, Defile the house, and fill the erant ante faciem domus 9:7 et dixit ad eos contaminate do-

courts with the slain: go ye forth. And they went forth, and slew in the city.

mum et implete atria interfectis egredimini et egressi sunt et

9:8 And it came to pass, while they were slaying them, and percutiebant eos qui erant in ciuitate 9:8 et caede conpleta

I was left, that I fell upon my face, and cried, and said, Ah Lord GOD! wilt thou destroy all the residue of Israel in thy pouring out of thy fury upon Jerusalem?

Prophetia Ezechielis

remansi ego ruique super faciem meam et clamans aio heu heu heu Domine Deus ergone disperdes omnes reliquias Israhel effundens furorem tuum super Hierusalem 9:9 et dixit 9:9 Then said he unto me, The iniquity of the house of

Israel and Judah is exceeding great, and the land is full of blood, and the city full of perverseness: for they say, The

ad me iniquitas domus Israhel et Iuda magna est nimis ualde LORD hath forsaken the earth, and the LORD seeth not. et repleta est terra sanguinibus et ciuitas repleta est auersione dixerunt enim dereliquit Dominus terram et Dominus non uidet 9:10 igitur et meus non parcet oculus neque miserebor 9:10 And as for me also, mine eye shall not spare, neither

will I have pity, but I will recompense their way upon their head.

uiam eorum super caput eorum reddam 9:11 et ecce uir qui

9:11 And, behold, the man clothed with linen, which had the inkhorn by his side, reported the matter, saying, I have done as thou hast commanded me.

indutus erat lineis qui habebat atramentarium in dorso suo respondit uerbum dicens feci sicut praecepisti mihi

10:1 et uidi et ecce in firmamento quod erat super caput 10:1 Then I looked, and, behold, in the firmament that was above the head of the cherubims there appeared over them as it were a sapphire stone, as the appearance of the like-

cherubin quasi lapis sapphyrus quasi species similitudinis ness of a throne. solii apparuit super ea 10:2 et dixit ad uirum qui indutus 10:2 And he spake unto the man clothed with linen, and

said, Go in between the wheels, even under the cherub, and fill thine hand with coals of fire from between the cherubims, and scatter them over the city. And he went in in my sight.

erat lineis et ait ingredere in medio rotarum quae sunt subtus cherub et imple manum tuam prunis ignis quae sunt inter cherubin et effunde super ciuitatem ingressusque est in conspectu meo 10:3 cherubin autem stabant a dextris domus 10:3 Now the cherubims stood on the right side of the

house, when the man went in; and the cloud filled the inner court.

cum ingrederetur uir et nubes impleuit atrium interius 10:4

10:4 Then the glory of the LORD went up from the cherub, and stood over the threshold of the house; and the house was filled with the cloud, and the court was full of the brightness of the LORD’s glory.

et eleuata est gloria Domini desuper cherub ad limen domus et repleta est domus nube et atrium repletum est splendore gloriae Domini 10:5 et sonitus alarum cherubin audiebatur 10:5 And the sound of the cherubims’ wings was heard

even to the outer court, as the voice of the Almighty God when he speaketh.

usque ad atrium exterius quasi uox Dei omnipotentis loquentis 10:6 cumque praecepisset uiro qui indutus erat lineis di- 10:6 And it came to pass, that when he had commanded

the man clothed with linen, saying, Take fire from between the wheels, from between the cherubims; then he went in,

cens sume ignem de medio rotarum quae sunt inter cheru- and stood beside the wheels. bin ingressus ille stetit iuxta rotam 10:7 et extendit cherub 10:7 And one cherub stretched forth his hand from be-

tween the cherubims unto the fire that was between the cherubims, and took thereof, and put it into the hands of him that was clothed with linen: who took it, and went out.

manum de medio cherubin ad ignem qui erat inter cherubin et sumpsit et dedit in manus eius qui indutus erat lineis qui accipiens egressus est 10:8 et apparuit in cherubin si- 10:8 And there appeared in the cherubims the form of a

man’s hand under their wings.

militudo manus hominis subtus pinnas eorum 10:9 et uidi 10:9 And when I looked, behold the four wheels by the

cherubims, one wheel by one cherub, and another wheel by another cherub: and the appearance of the wheels was

et ecce quattuor rotae iuxta cherubin rota una iuxta cherub as the colour of a beryl stone. unum et rota alia iuxta cherub unum species autem erat rotarum quasi uisio lapidis chrysoliti 10:10 et aspectus earum 10:10 And as for their appearances, they four had one like-

ness, as if a wheel had been in the midst of a wheel.

similitudo una quattuor quasi sit rota in medio rotae 10:11 10:11 When they went, they went upon their four sides;

they turned not as they went, but to the place whither the head looked they followed it; they turned not as they went.

Biblia Sacra Vulgata

cumque ambularent in quattuor partes gradiebantur non reuertebantur ambulantes sed ad locum ad quem ire declinabat quae prima erat sequebantur et ceterae nec conuertebantur

10:12 And their whole body, and their backs, and their 10:12 et omne corpus earum et colla et manus et pinnae et

hands, and their wings, and the wheels, were full of eyes round about, even the wheels that they four had.

circuli plena erant oculis in circuitu quattuor rotarum 10:13

10:13 As for the wheels, it was cried unto them in my hearing, O wheel.

10:14 And every one had four faces: the first face was the et rotas istas uocauit uolubiles audiente me 10:14 quattuor

face of a cherub, and the second face was the face of a man, and the third the face of a lion, and the fourth the

face of an eagle. autem facies habebat unum facies una facies cherub et fa-

cies secunda facies hominis et in tertio facies leonis et in

10:15 And the cherubims were lifted up. This is the living quarto facies aquilae 10:15 et eleuata sunt cherubin ipsum

creature that I saw by the river of Chebar.

10:16 And when the cherubims went, the wheels went by est animal quod uideram iuxta flumen Chobar 10:16 cum-

them: and when the cherubims lifted up their wings to mount up from the earth, the same wheels also turned not

from beside them. que ambularent cherubin ibant pariter et rotae iuxta ea et

cum leuarent cherubin alas suas ut exaltarentur de terra non

10:17 When they stood, these stood; and when they were residebant rotae sed et ipsae iuxta erant 10:17 stantibus illis

lifted up, these lifted up themselves also: for the spirit of the living creature was in them.

stabant et cum eleuatis eleuabantur spiritus enim uitae erat in

10:18 Then the glory of the LORD departed from off the eis 10:18 et egressa est gloria Domini a limine templi et ste-

threshold of the house, and stood over the cherubims.

10:19 And the cherubims lifted up their wings, and tit super cherubin 10:19 et eleuantia cherubin alas suas ex-

mounted up from the earth in my sight: when they went out, the wheels also were beside them, and every one stood at the door of the east gate of the LORD’s house; and the glory of the God of Israel was over them above.

altata sunt a terra coram me et illis egredientibus rotae quoque subsecutae sunt et stetit in introitu portae domus Domini

10:20 This is the living creature that I saw under the God orientalis et gloria Dei Israhel erat super ea 10:20 ipsum est

of Israel by the river of Chebar; and I knew that they were the cherubims.

animal quod uidi subter Deum Israhel iuxta fluuium Chobar

10:21 Every one had four faces apiece, and every one four et intellexi quia cherubin essent 10:21 quattuor per quattuor

wings; and the likeness of the hands of a man was under their wings.

uultus uni et quattuor alae uni et similitudo manus hominis

10:22 And the likeness of their faces was the same faces sub alis eorum 10:22 et similitudo uultuum eorum ipsi uul-

which I saw by the river of Chebar, their appearances and themselves: they went every one straight forward.

tus quos uideram iuxta fluuium Chobar et intuitus eorum et impetus singulorum ante faciem suam ingredi

11:1 Moreover the spirit lifted me up, and brought me unto 11:1 et eleuauit me spiritus et introduxit me ad portam do-

the east gate of the LORD’s house, which looketh eastward: and behold at the door of the gate five and twenty men; among whom I saw Jaazaniah the son of Azur, and Pelatiah the son of Benaiah, princes of the people.

mus Domini orientalem quae respicit solis ortum et ecce in introitu portae uiginti quinque uiri et uidi in medio eorum Hiezoniam filium Azur et Pheltiam filium Banaiae principes

11:2 Then said he unto me, Son of man, these are the men populi 11:2 dixitque ad me fili hominis hii uiri qui cogi-

that devise mischief, and give wicked counsel in this city:

tant iniquitatem et tractant consilium pessimum in urbe ista

11:3 Which say, It is not near; let us build houses: this city 11:3 dicentes nonne dudum aedificatae sunt domus haec est is the caldron, and we be the flesh.

11:4 Therefore prophesy against them, prophesy, O son of lebes nos autem carnes 11:4 idcirco uaticinare de eis uatici-

man.

Prophetia Ezechielis

nare fili hominis 11:5 et inruit in me spiritus Domini et dixit 11:5 And the Spirit of the LORD fell upon me, and said

unto me, Speak; Thus saith the LORD; Thus have ye said, O house of Israel: for I know the things that come into

ad me loquere haec dicit Dominus sic locuti estis domus Is- your mind, every one of them. rahel et cogitationes cordis uestri ego noui 11:6 plurimos 11:6 Ye have multiplied your slain in this city, and ye have

filled the streets thereof with the slain.

occidistis in urbe hac et implestis uias eius interfectis 11:7 11:7 Therefore thus saith the Lord GOD; Your slain whom

ye have laid in the midst of it, they are the flesh, and this city is the caldron: but I will bring you forth out of the

propterea haec dicit Dominus Deus interfecti uestri quos po- midst of it. suistis in medio eius hii sunt carnes et haec est lebes et educam uos de medio eius 11:8 gladium metuistis et gladium 11:8 Ye have feared the sword; and I will bring a sword

upon you, saith the Lord GOD.

inducam super uos ait Dominus Deus 11:9 et eiciam uos de 11:9 And I will bring you out of the midst thereof, and

deliver you into the hands of strangers, and will execute judgments among you.

medio eius daboque uos in manu hostium et faciam in uobis iudicia 11:10 gladio cadetis in finibus Israhel iudicabo 11:10 Ye shall fall by the sword; I will judge you in the

border of Israel; and ye shall know that I am the LORD.

uos et scietis quia ego Dominus 11:11 haec non erit uobis 11:11 This city shall not be your caldron, neither shall ye be the flesh in the midst thereof; but I will judge you in the border of Israel:

in lebetem et uos non eritis in medio eius in carnes in finibus Israhel iudicabo uos 11:12 et scietis quia ego Dominus 11:12 And ye shall know that I am the LORD: for ye have

not walked in my statutes, neither executed my judgments, but have done after the manners of the heathen that are

qui in praeceptis meis non ambulastis et iudicia mea non fe- round about you. cistis sed iuxta iudicia gentium quae in circuitu uestro sunt estis operati 11:13 et factum est cum prophetarem Pheltias 11:13 And it came to pass, when I prophesied, that Pela-

tiah the son of Benaiah died. Then fell I down upon my face, and cried with a loud voice, and said, Ah Lord GOD!

filius Banaiae mortuus est et cecidi in faciem meam clamans wilt thou make a full end of the remnant of Israel? uoce magna et dixi heu heu heu Domine Deus consummationem tu facis reliquiarum Israhel 11:14 et factum est uer- 11:14 Again the word of the LORD came unto me, saying, bum Domini ad me dicens 11:15 fili hominis fratres tui fra- 11:15 Son of man, thy brethren, even thy brethren, the

men of thy kindred, and all the house of Israel wholly, are they unto whom the inhabitants of Jerusalem have said, Get you far from the LORD: unto us is this land given in possession.

tres tui uiri propinqui tui et omnis domus Israhel uniuersi quibus dixerunt habitatores Hierusalem longe recedite a Domino nobis data est terra in possessionem 11:16 propterea 11:16 Therefore say, Thus saith the Lord GOD; Although

I have cast them far off among the heathen, and although I have scattered them among the countries, yet will I be to them as a little sanctuary in the countries where they shall come.

haec dicit Dominus Deus quia longe feci eos in gentibus et quia dispersi eos in terris ero eis in sanctificationem modicam in terris ad quas uenerunt 11:17 propterea loquere haec 11:17 Therefore say, Thus saith the Lord GOD; I will even

gather you from the people, and assemble you out of the countries where ye have been scattered, and I will give you

dicit Dominus Deus congregabo uos de populis et adunabo the land of Israel. de terris in quibus dispersi estis daboque uobis humum Israhel 11:18 et ingredientur illuc et auferent omnes offensio- 11:18 And they shall come thither, and they shall take

away all the detestable things thereof and all the abominations thereof from thence.

nes cunctasque abominationes eius de illa 11:19 et dabo eis

11:19 And I will give them one heart, and I will put a new spirit within you; and I will take the stony heart out of their flesh, and will give them an heart of flesh:

cor unum et spiritum nouum tribuam in uisceribus eorum et auferam cor lapideum de carne eorum et dabo eis cor carneum 11:20 ut in praeceptis meis ambulent et iudicia mea 11:20 That they may walk in my statutes, and keep mine

ordinances, and do them: and they shall be my people, and I will be their God.

custodiant faciantque ea et sint mihi in populum et ego sim

11:21 But as for them whose heart walketh after the heart eis in Deum 11:21 quorum cor post offendicula et abomiof their detestable things and their abominations, I will recompense their way upon their own heads, saith the Lord

GOD. nationes suas ambulat horum uiam in capite suo ponam dicit

11:22 Then did the cherubims lift up their wings, and the Dominus Deus 11:22 et eleuauerunt cherubin alas suas et wheels beside them; and the glory of the God of Israel was over them above.

rotae cum eis et gloria Dei Israhel erat super ea 11:23 et as-

11:23 And the glory of the LORD went up from the midst of the city, and stood upon the mountain which is on the east side of the city.

cendit gloria Domini de medio ciuitatis stetitque super mon-

11:24 Afterwards the spirit took me up, and brought me in tem qui est ad orientem urbis 11:24 et spiritus leuauit me

a vision by the Spirit of God into Chaldea, to them of the captivity. So the vision that I had seen went up from me.

adduxitque in Chaldeam ad transmigrationem in uisione in

11:25 Then I spake unto them of the captivity all the things spiritu Dei et sublata est a me uisio quam uideram 11:25 et that the LORD had shewed me.

locutus sum ad transmigrationem omnia uerba Domini quae ostenderat mihi

12:1 The word of the LORD also came unto me, saying, 12:1 et factus est sermo Domini ad me dicens 12:2 fili

12:2 Son of man, thou dwellest in the midst of a rebellious house, which have eyes to see, and see not; they have ears to hear, and hear not: for they are a rebellious house.

hominis in medio domus exasperantis tu habitas qui oculos habent ad uidendum et non uident et aures ad audiendum

12:3 Therefore, thou son of man, prepare thee stuff for re- et non audiunt quia domus exasperans est 12:3 tu ergo fili moving, and remove by day in their sight; and thou shalt remove from thy place to another place in their sight: it may be they will consider, though they be a rebellious house.

hominis fac tibi uasa transmigrationis et transmigrabis per diem coram eis transmigrabis autem de loco tuo ad locum alterum in conspectu eorum si forte aspiciant quia domus ex-

12:4 Then shalt thou bring forth thy stuff by day in their asperans est 12:4 et efferes foras uasa tua quasi uasa trans-

sight, as stuff for removing: and thou shalt go forth at even in their sight, as they that go forth into captivity.

migrantis per diem in conspectu eorum tu autem egredieris

12:5 Dig thou through the wall in their sight, and carry out uespere coram eis sicut egreditur migrans 12:5 ante oculos

thereby.

12:6 In their sight shalt thou bear it upon thy shoulders, eorum perfodi tibi parietem et egredieris per eum 12:6 in and carry it forth in the twilight: thou shalt cover thy face, that thou see not the ground: for I have set thee for a sign

unto the house of Israel. conspectu eorum in umeris portaberis in caligine effereris faciem tuam uelabis et non uidebis terram quia portentum dedi

12:7 And I did so as I was commanded: I brought forth my te domui Israhel 12:7 feci ergo sicut praeceperat mihi uasa

stuff by day, as stuff for captivity, and in the even I digged through the wall with mine hand; I brought it forth in the

twilight, and I bare it upon my shoulder in their sight. mea protuli quasi uasa transmigrantis per diem et uespere perfodi mihi parietem manu in caligine egressus sum et in

12:8 And in the morning came the word of the LORD unto umeris portatus in conspectu eorum 12:8 et factus est sermo

me, saying,

12:9 Son of man, hath not the house of Israel, the rebel- Domini ad me mane dicens 12:9 fili hominis numquid non lious house, said unto thee, What doest thou?

dixerunt ad te domus Israhel domus exasperans quid tu fa-

12:10 Say thou unto them, Thus saith the Lord GOD; cis 12:10 dic ad eos haec dicit Dominus Deus super ducem

This burden concerneth the prince in Jerusalem, and all the house of Israel that are among them.

onus istud qui est in Hierusalem et super omnem domum

12:11 Say, I am your sign: like as I have done, so shall it be Israhel quae est in medio eorum 12:11 dic ego portentum done unto them: they shall remove and go into captivity.

Prophetia Ezechielis

uestrum quomodo feci sic fiet illis in transmigrationem et captiuitatem ibunt 12:12 et dux qui est in medio eorum in 12:12 And the prince that is among them shall bear upon

his shoulder in the twilight, and shall go forth: they shall dig through the wall to carry out thereby: he shall cover

umeris portabitur in caligine egredietur parietem perfodient his face, that he see not the ground with his eyes. ut educant eum facies eius operietur ut non uideat oculo terram 12:13 et extendam rete meum super illum et capietur in 12:13 My net also will I spread upon him, and he shall be

taken in my snare: and I will bring him to Babylon to the land of the Chaldeans; yet shall he not see it, though he

sagena mea et adducam eum in Babylonem in terram Chal- shall die there. deorum et ipsam non uidebit ibique morietur 12:14 et om- 12:14 And I will scatter toward every wind all that are

about him to help him, and all his bands; and I will draw out the sword after them.

nes qui circa eum sunt praesidium eius et agmina eius dispergam in omnem uentum et gladium euaginabo post eos 12:15 et scient quia ego Dominus quando dispersero illos in 12:15 And they shall know that I am the LORD, when I

shall scatter them among the nations, and disperse them in the countries.

gentibus et disseminauero eos in terris 12:16 et relinquam

12:16 But I will leave a few men of them from the sword, from the famine, and from the pestilence; that they may declare all their abominations among the heathen whither they come; and they shall know that I am the LORD.

ex eis uiros paucos a gladio et fame et pestilentia ut narrent omnia scelera eorum in gentibus ad quas ingredientur et scient quia ego Dominus 12:17 et factus est sermo Domini 12:17 Moreover the word of the LORD came to me, say-

ing,

ad me dicens 12:18 fili hominis panem tuum in conturba- 12:18 Son of man, eat thy bread with quaking, and drink

thy water with trembling and with carefulness;

tione comede sed et aquam tuam in festinatione et maerore bibe 12:19 et dices ad populum terrae haec dicit Dominus 12:19 And say unto the people of the land, Thus saith the

Lord GOD of the inhabitants of Jerusalem, and of the land of Israel; They shall eat their bread with carefulness, and drink their water with astonishment, that her land may be desolate from all that is therein, because of the violence of all them that dwell therein.

Deus ad eos qui habitant in Hierusalem in terra Israhel panem suum in sollicitudine comedent et aquam suam in desolatione bibent ut desoletur terra a multitudine sua prop-

ter iniquitatem omnium qui habitant in ea 12:20 et ciuita- 12:20 And the cities that are inhabited shall be laid waste, and the land shall be desolate; and ye shall know that I am the LORD.

tes quae nunc habitantur desolatae erunt terraque deserta et scietis quia ego Dominus 12:21 et factus est sermo Domini 12:21 And the word of the LORD came unto me, saying, ad me dicens 12:22 fili hominis quod est prouerbium istud 12:22 Son of man, what is that proverb that ye have in the

land of Israel, saying, The days are prolonged, and every vision faileth?

uobis in terra Israhel dicentium in longum differentur dies et peribit omnis uisio 12:23 ideo dic ad eos haec dicit Do- 12:23 Tell them therefore, Thus saith the Lord GOD; I will

make this proverb to cease, and they shall no more use it as a proverb in Israel; but say unto them, The days are at

minus Deus quiescere faciam prouerbium istud neque uulgo hand, and the effect of every vision. dicetur ultra in Israhel et loquere ad eos quod adpropinquauerint dies et sermo omnis uisionis 12:24 non enim erit ultra 12:24 For there shall be no more any vain vision nor flat-

tering divination within the house of Israel.

omnis uisio cassa neque diuinatio ambigua in medio filiorum Israhel 12:25 quia ego Dominus loquar quodcumque 12:25 For I am the LORD: I will speak, and the word that

I shall speak shall come to pass; it shall be no more prolonged: for in your days, O rebellious house, will I say the

locutus fuero uerbum et fiet non prolongabitur amplius sed word, and will perform it, saith the Lord GOD. in diebus uestris domus exasperans loquar uerbum et faciam

Biblia Sacra Vulgata

12:26 Again the word of the LORD came to me, saying. illud dicit Dominus Deus 12:26 et factus est sermo Domini

12:27 Son of man, behold, they of the house of Israel say, ad me dicens 12:27 fili hominis ecce domus Israhel dicen-The vision that he seeth is for many days to come, and he prophesieth of the times that are far off.

tium uisio quam hic uidet in dies multos et in tempora longa

12:28 Therefore say unto them, Thus saith the Lord GOD; iste prophetat 12:28 propterea dic ad eos haec dicit Domi-

There shall none of my words be prolonged any more, but the word which I have spoken shall be done, saith the Lord

GOD. nus Deus non prolongabitur ultra omnis sermo meus uerbum quod locutus fuero conplebitur dicit Dominus Deus

13:1 And the word of the LORD came unto me, saying, 13:1 et factus est sermo Domini ad me dicens 13:2 fili

13:2 Son of man, prophesy against the prophets of Israel that prophesy, and say thou unto them that prophesy out of their own hearts, Hear ye the word of the LORD;

hominis uaticinare ad prophetas Israhel qui prophetant et dices prophetantibus de corde suo audite uerbum Domini

13:3 Thus saith the Lord GOD; Woe unto the foolish 13:3 haec dicit Dominus Deus uae prophetis insipientibus qui

prophets, that follow their own spirit, and have seen nothing!

sequuntur spiritum suum et nihil uident 13:4 quasi uulpes

13:4 O Israel, thy prophets are like the foxes in the deserts.

13:5 Ye have not gone up into the gaps, neither made up in desertis prophetae tui Israhel erant 13:5 non ascendis-

the hedge for the house of Israel to stand in the battle in the day of the LORD.

tis ex aduerso neque opposuistis murum pro domo Israhel

13:6 They have seen vanity and lying divination, saying, ut staretis in proelio in die Domini 13:6 uident uana et di-

The LORD saith: and the LORD hath not sent them: and they have made others to hope that they would confirm the

word. uinant mendacium dicentes ait Dominus cum Dominus non 13:7 Have ye not seen a vain vision, and have ye not spo- miserit eos et perseuerauerunt confirmare sermonem 13:7

ken a lying divination, whereas ye say, The LORD saith it; albeit I have not spoken?

numquid non uisionem cassam uidistis et diuinationem mendacem locuti estis et dicitis ait Dominus cum ego non sim

13:8 Therefore thus saith the Lord GOD; Because ye locutus 13:8 propterea haec dicit Dominus Deus quia lo-

have spoken vanity, and seen lies, therefore, behold, I am against you, saith the Lord GOD.

cuti estis uana et uidistis mendacium ideo ecce ego ad uos

13:9 And mine hand shall be upon the prophets that see ait Dominus Deus 13:9 et erit manus mea super prophetas

vanity, and that divine lies: they shall not be in the assembly of my people, neither shall they be written in the writing of the house of Israel, neither shall they enter into the land of Israel; and ye shall know that I am the Lord GOD.

qui uident uana et diuinant mendacium in concilio populi mei non erunt et in scriptura domus Israhel non scribentur nec in terra Israhel ingredientur et scietis quia ego Domi-

13:10 Because, even because they have seduced my peo- nus Deus 13:10 eo quod deceperint populum meum dicenple, saying, Peace; and there was no peace; and one built up a wall, and, lo, others daubed it with untempered

morter: tes pax et non est pax et ipse aedificabat parietem illi au-13:11 Say unto them which daub it with untempered tem liniebant eum luto absque paleis 13:11 dic ad eos qui

morter, that it shall fall: there shall be an overflowing shower; and ye, O great hailstones, shall fall; and a stormy

wind shall rend it. liniunt absque temperatura quod casurus sit erit enim im-

ber inundans et dabo lapides praegrandes desuper inruentes

13:12 Lo, when the wall is fallen, shall it not be said unto et uentum procellae dissipantem 13:12 siquidem ecce ceyou, Where is the daubing wherewith ye have daubed it?

cidit paries numquid non dicetur uobis ubi est litura quam

13:13 Therefore thus saith the Lord GOD; I will even rend leuistis 13:13 propterea haec dicit Dominus Deus et erum-

it with a stormy wind in my fury; and there shall be an overflowing shower in mine anger, and great hailstones in

my fury to consume it. pere faciam spiritum tempestatum in indignatione mea et im-

Prophetia Ezechielis

ber inundans in furore meo erit et lapides grandes in ira in consummationem 13:14 et destruam parietem quem leuistis 13:14 So will I break down the wall that ye have daubed

with untempered morter, and bring it down to the ground, so that the foundation thereof shall be discovered, and it shall fall, and ye shall be consumed in the midst thereof: and ye shall know that I am the LORD.

absque temperamento et adaequabo eum terrae et reuelabitur fundamentum eius et cadet et consumetur in medio eius et scietis quia ego sum Dominus 13:15 et conplebo indig- 13:15 Thus will I accomplish my wrath upon the wall, and

upon them that have daubed it with untempered morter, and will say unto you, The wall is no more, neither they

nationem meam in parietem et in his qui linunt eum absque that daubed it; temperamento dicamque uobis non est paries et non sunt qui linunt eum 13:16 prophetae Israhel qui prophetant ad Hie- 13:16 To wit, the prophets of Israel which prophesy con-

cerning Jerusalem, and which see visions of peace for her, and there is no peace, saith the Lord GOD.

rusalem et uident ei uisionem pacis et non est pax ait Domi-

nus Deus 13:17 et tu fili hominis pone faciem tuam contra 13:17 Likewise, thou son of man, set thy face against the daughters of thy people, which prophesy out of their own heart; and prophesy thou against them,

filias populi tui quae prophetant de corde suo et uaticinare super eas 13:18 et dic haec ait Dominus Deus uae quae con- 13:18 And say, Thus saith the Lord GOD; Woe to the

women that sew pillows to all armholes, and make kerchiefs upon the head of every stature to hunt souls! Will ye hunt the souls of my people, and will ye save the souls alive that come unto you?

suunt puluillos sub omni cubito manus et faciunt ceruicalia sub capite uniuersae aetatis ad capiendas animas cum caperent animas populi mei uiuificabant animas eorum 13:19 et 13:19 And will ye pollute me among my people for hand-

fuls of barley and for pieces of bread, to slay the souls that should not die, and to save the souls alive that should not

uiolabant me ad populum meum propter pugillum hordei et live, by your lying to my people that hear your lies? fragmen panis ut interficerent animas quae non moriuntur et uiuificarent animas quae non uiuunt mentientes populo meo credenti mendaciis 13:20 propter hoc haec dicit Dominus 13:20 Wherefore thus saith the Lord GOD; Behold, I am

against your pillows, wherewith ye there hunt the souls to make them fly, and I will tear them from your arms, and will let the souls go, even the souls that ye hunt to make them fly.

Deus ecce ego ad puluillos uestros quibus uos capitis animas uolantes et disrumpam eos de brachiis uestris et dimit-

tam animas quas uos capitis animas ad uolandum 13:21 et 13:21 Your kerchiefs also will I tear, and deliver my people out of your hand, and they shall be no more in your hand to be hunted; and ye shall know that I am the LORD.

disrumpam ceruicalia uestra et liberabo populum meum de manu uestra neque erunt ultra in manibus uestris ad praedan-

dum et scietis quia ego Dominus 13:22 pro eo quod maerere 13:22 Because with lies ye have made the heart of the righteous sad, whom I have not made sad; and strengthened the hands of the wicked, that he should not return from his

fecistis cor iusti mendaciter quem ego non contristaui et con- wicked way, by promising him life: fortastis manus impii ut non reuerteretur a uia sua mala et uiueret 13:23 propterea uana non uidebitis et diuinationes 13:23 Therefore ye shall see no more vanity, nor divine

divinations: for I will deliver my people out of your hand: and ye shall know that I am the LORD.

non diuinabitis amplius et eruam populum meum de manu uestra et scietis quoniam ego Dominus

14:1 et uenerunt ad me uiri seniorum Israhel et sederunt 14:1 Then came certain of the elders of Israel unto me, and

sat before me.

coram me 14:2 et factus est sermo Domini ad me dicens 14:2 And the word of the LORD came unto me, saying, 14:3 fili hominis uiri isti posuerunt inmunditias suas in cor- 14:3 Son of man, these men have set up their idols in their

heart, and put the stumblingblock of their iniquity before their face: should I be enquired of at all by them?

Biblia Sacra Vulgata

dibus suis et scandalum iniquitatis suae statuerunt contra fa-

14:4 Therefore speak unto them, and say unto them, Thus ciem suam numquid interrogatus respondebo eis 14:4 pro-

saith the Lord GOD; Every man of the house of Israel that setteth up his idols in his heart, and putteth the stumblingblock of his iniquity before his face, and cometh to the prophet; I the LORD will answer him that cometh according to the multitude of his idols;

pter hoc loquere eis et dices ad eos haec dicit Dominus Deus homo homo de domo Israhel qui posuerit inmunditias suas in corde suo et scandalum iniquitatis suae statuerit contra faciem suam et uenerit ad prophetam interrogans per eum me ego Dominus respondebo ei in multitudine inmunditiarum

14:5 That I may take the house of Israel in their own heart, suarum 14:5 ut capiatur domus Israhel in corde suo quo rebecause they are all estranged from me through their idols. 14:6 Therefore say unto the house of Israel, Thus saith the cesserunt a me in cunctis idolis suis 14:6 propterea dic ad Lord GOD; Repent, and turn yourselves from your idols; and turn away your faces from all your abominations.

domum Israhel haec dicit Dominus Deus conuertimini et recedite ab idolis uestris et ab uniuersis contaminationibus ue-

14:7 For every one of the house of Israel, or of the stranger stris auertite facies uestras 14:7 quia homo homo de domo that sojourneth in Israel, which separateth himself from me, and setteth up his idols in his heart, and putteth the stumblingblock of his iniquity before his face, and cometh to a prophet to enquire of him concerning me; I the LORD will answer him by myself:

Israhel et de proselytis quicumque aduena fuerit in Israhel si alienatus fuerit a me et posuerit idola sua in corde suo et scandalum iniquitatis suae statuerit contra faciem suam et uenerit ad prophetam ut interroget per eum me ego Do-

14:8 And I will set my face against that man, and will minus respondebo ei per me 14:8 et ponam faciem meam make him a sign and a proverb, and I will cut him off from the midst of my people; and ye shall know that I am the

LORD. super hominem illum et faciam eum in exemplum et in pro-

uerbium et disperdam eum de medio populi mei et scietis

14:9 And if the prophet be deceived when he hath spoken quia ego Dominus 14:9 et propheta cum errauerit et locu-

a thing, I the LORD have deceived that prophet, and I will stretch out my hand upon him, and will destroy him from

the midst of my people Israel. tus fuerit uerbum ego Dominus decepi prophetam illum et

extendam manum meam super eum et delebo eum de me-

14:10 And they shall bear the punishment of their iniq- dio populi mei Israhel 14:10 et portabunt iniquitatem suam

uity: the punishment of the prophet shall be even as the punishment of him that seeketh unto him;

iuxta iniquitatem interrogantis sic iniquitas prophetae erit

14:11 That the house of Israel may go no more astray from 14:11 ut non erret ultra domus Israhel a me neque polluatur

me, neither be polluted any more with all their transgressions; but that they may be my people, and I may be their

God, saith the Lord GOD. in uniuersis praeuaricationibus suis sed sit mihi in populum 14:12 The word of the LORD came again to me, saying, et ego sim eis in Deum ait Dominus exercituum 14:12 et 14:13 Son of man, when the land sinneth against me by factus est sermo Domini ad me dicens 14:13 fili hominis

trespassing grievously, then will I stretch out mine hand upon it, and will break the staff of the bread thereof, and will send famine upon it, and will cut off man and beast from it:

terra cum peccauerit mihi ut praeuaricetur praeuaricans extendam manum meam super eam et conteram uirgam panis eius et inmittam in eam famem et interficiam de ea homi-

14:14 Though these three men, Noah, Daniel, and Job, nem et iumentum 14:14 et si fuerint tres uiri isti in medio

were in it, they should deliver but their own souls by their righteousness, saith the Lord GOD.

eius Noe Danihel et Iob ipsi iustitia sua liberabunt animas

Prophetia Ezechielis

suas ait Dominus exercituum 14:15 quod si et bestias pes- 14:15 If I cause noisome beasts to pass through the land,

and they spoil it, so that it be desolate, that no man may pass through because of the beasts:

simas induxero super terram ut uastem eam et fuerit inuia eo quod non sit pertransiens propter bestias 14:16 tres uiri 14:16 Though these three men were in it, as I live, saith the

Lord GOD, they shall deliver neither sons nor daughters; they only shall be delivered, but the land shall be desolate.

isti qui fuerint in ea uiuo ego dicit Dominus Deus quia nec filios nec filias liberabunt sed ipsi soli liberabuntur terra autem desolabitur 14:17 uel si gladium induxero super terram 14:17 Or if I bring a sword upon that land, and say, Sword,

go through the land; so that I cut off man and beast from it:

illam et dixero gladio transi per terram et interfecero de ea

hominem et iumentum 14:18 et tres uiri isti fuerint in me- 14:18 Though these three men were in it, as I live, saith the Lord GOD, they shall deliver neither sons nor daughters, but they only shall be delivered themselves.

dio eius uiuo ego dicit Dominus Deus non liberabunt filios neque filias sed ipsi soli liberabuntur 14:19 si autem et pes- 14:19 Or if I send a pestilence into that land, and pour out

my fury upon it in blood, to cut off from it man and beast:

tilentiam inmisero super terram illam et effudero indignationem meam super eam in sanguine ut auferam ex ea hominem et iumentum 14:20 et Noe et Danihel et Iob fuerint in medio 14:20 Though Noah, Daniel, and Job were in it, as I live,

saith the Lord GOD, they shall deliver neither son nor daughter; they shall but deliver their own souls by their

eius uiuo ego dicit Dominus Deus quia filium et filiam non righteousness. liberabunt sed ipsi iustitia sua liberabunt animas suas 14:21 14:21 For thus saith the Lord GOD; How much more when

I send my four sore judgments upon Jerusalem, the sword, and the famine, and the noisome beast, and the pestilence,

quoniam haec dicit Dominus Deus quod si et quattuor iudi- to cut off from it man and beast? cia mea pessima gladium et famem et bestias malas et pestilentiam misero in Hierusalem ut interficiam de ea hominem

et pecus 14:22 tamen relinquetur in ea saluatio educentium 14:22 Yet, behold, therein shall be left a remnant that shall be brought forth, both sons and daughters: behold, they shall come forth unto you, and ye shall see their way and their doings: and ye shall be comforted concerning the evil that I have brought upon Jerusalem, even concerning all that I have brought upon it.

filios et filias ecce ipsi egredientur ad uos et uidebitis uiam eorum et adinuentiones eorum et consolabimini super malo quod induxi in Hierusalem in omnibus quae inportaui super eam 14:23 et consolabuntur uos cum uideritis uiam eorum 14:23 And they shall comfort you, when ye see their ways

and their doings: and ye shall know that I have not done without cause all that I have done in it, saith the Lord

et adinuentiones eorum et cognoscetis quod non frustra fe- GOD. cerim omnia quae feci in ea ait Dominus Deus

15:1 et factus est sermo Domini ad me dicens 15:2 fili ho- 15:1 And the word of the LORD came unto me, saying,

15:2 Son of man, what is the vine tree more than any tree, or than a branch which is among the trees of the forest?

minis quid fiet ligno uitis ex omnibus lignis nemorum quae sunt inter ligna siluarum 15:3 numquid tolletur de ea lig- 15:3 Shall wood be taken thereof to do any work? or will

men take a pin of it to hang any vessel thereon?

num ut fiat opus aut fabricabitur de ea paxillus ut dependeat

in eo quodcumque uas 15:4 ecce igni datum est in escam 15:4 Behold, it is cast into the fire for fuel; the fire devoureth both the ends of it, and the midst of it is burned. Is it meet for any work?

utramque partem eius consumpsit ignis et medietas eius redacta est in fauillam numquid utile erit ad opus 15:5 etiam 15:5 Behold, when it was whole, it was meet for no work:

how much less shall it be meet yet for any work, when the fire hath devoured it, and it is burned?

cum esset integrum non erat aptum ad opus quanto magis

Biblia Sacra Vulgata

cum ignis illud deuorauerit et conbuserit nihil ex eo fiet ope-

15:6 Therefore thus saith the Lord GOD; As the vine tree ris 15:6 propterea haec dicit Dominus Deus quomodo lig-

among the trees of the forest, which I have given to the fire for fuel, so will I give the inhabitants of Jerusalem.

num uitis inter ligna siluarum quod dedi igni ad deuorandum

15:7 And I will set my face against them; they shall go out sic tradidi habitatores Hierusalem 15:7 et ponam faciem

from one fire, and another fire shall devour them; and ye shall know that I am the LORD, when I set my face against

them. meam in eos de igne egredientur et ignis consumet eos et

scietis quia ego Dominus cum posuero faciem meam in eos

15:8 And I will make the land desolate, because they have 15:8 et dedero terram inuiam et desolatam eo quod praeuari-

committed a trespass, saith the Lord GOD.

catores extiterint dicit Dominus Deus

16:1 Again the word of the LORD came unto me, saying, 16:1 et factus est sermo Domini ad me dicens 16:2 fili 16:2 Son of man, cause Jerusalem to know her abominations,

hominis notas fac Hierusalem abominationes suas 16:3 et

16:3 And say, Thus saith the Lord GOD unto Jerusalem; Thy birth and thy nativity is of the land of Canaan; thy

father was an Amorite, and thy mother an Hittite. dices haec dicit Dominus Deus Hierusalem radix tua et ge-

neratio tua de terra chananea pater tuus Amorreus et mater

16:4 And as for thy nativity, in the day thou wast born thy tua Cetthea 16:4 et quando nata es in die ortus tui non est navel was not cut, neither wast thou washed in water to supple thee; thou wast not salted at all, nor swaddled at

all. praecisus umbilicus tuus et in aqua non es lota in salutem 16:5 None eye pitied thee, to do any of these unto thee, to nec sale salita nec inuoluta pannis 16:5 non pepercit super

have compassion upon thee; but thou wast cast out in the open field, to the lothing of thy person, in the day that thou

wast born. te oculus ut facerem tibi unum de his miseratus tui sed pro-

iecta es super faciem terrae in abiectione animae tuae in die

16:6 And when I passed by thee, and saw thee polluted in qua nata es 16:6 transiens autem per te uidi te conculcari in

thine own blood, I said unto thee when thou wast in thy blood, Live; yea, I said unto thee when thou wast in thy

blood, Live. sanguine tuo et dixi tibi cum esses in sanguine tuo uiue dixi 16:7 I have caused thee to multiply as the bud of the field, inquam tibi in sanguine tuo uiue 16:7 multiplicatam quasi

and thou hast increased and waxen great, and thou art come to excellent ornaments: thy breasts are fashioned,

and thine hair is grown, whereas thou wast naked and bare. germen agri dedi te et multiplicata es et grandis effecta et ingressa es et peruenisti ad mundum muliebrem ubera tua intumuerunt et pilus tuus germinauit et eras nuda et confu-

16:8 Now when I passed by thee, and looked upon thee, sionis plena 16:8 et transiui per te et uidi te et ecce tempus

behold, thy time was the time of love; and I spread my skirt over thee, and covered thy nakedness: yea, I sware unto thee, and entered into a covenant with thee, saith the Lord GOD, and thou becamest mine.

tuum tempus amantium et expandi amictum meum super te et operui ignominiam tuam et iuraui tibi et ingressus sum

16:9 Then washed I thee with water; yea, I throughly pactum tecum ait Dominus Deus et facta es mihi 16:9 et

washed away thy blood from thee, and I anointed thee with oil.

laui te aqua et emundaui sanguinem tuum ex te et unxi te

16:10 I clothed thee also with broidered work, and shod oleo 16:10 et uestiui te discoloribus et calciaui te ianthino

thee with badgers’ skin, and I girded thee about with fine linen, and I covered thee with silk.

et cinxi te bysso et indui te subtilibus 16:11 et ornaui te or-

16:11 I decked thee also with ornaments, and I put bracelets upon thy hands, and a chain on thy neck.

namento et dedi armillas in manibus tuis et torquem circa

16:12 And I put a jewel on thy forehead, and earrings in collum tuum 16:12 et dedi inaurem super os tuum et circu-

thine ears, and a beautiful crown upon thine head.

16:13 Thus wast thou decked with gold and silver; and thy los auribus tuis et coronam decoris in capite tuo 16:13 et

raiment was of fine linen, and silk, and broidered work; thou didst eat fine flour, and honey, and oil: and thou wast exceeding beautiful, and thou didst prosper into a kingdom.

Prophetia Ezechielis

ornata es auro et argento et uestita es bysso et polymito et multicoloribus similam et mel et oleum comedisti et decora facta es uehementer nimis et profecisti in regnum 16:14 et 16:14 And thy renown went forth among the heathen for

thy beauty: for it was perfect through my comeliness, which I had put upon thee, saith the Lord GOD.

egressum est nomen tuum in gentes propter speciem tuam quia perfecta eras in decore meo quem posueram super te dicit Dominus Deus 16:15 et habens fiduciam in pulchritu- 16:15 But thou didst trust in thine own beauty, and

playedst the harlot because of thy renown, and pouredst out thy fornications on every one that passed by; his it

dine tua fornicata es in nomine tuo et exposuisti fornicatio- was. nem tuam omni transeunti ut eius fieres 16:16 et sumens de 16:16 And of thy garments thou didst take, and deckedst

thy high places with divers colours, and playedst the harlot thereupon: the like things shall not come, neither shall it

uestimentis meis fecisti tibi excelsa hinc inde consuta et for- be so. nicata es super eis sicut non est factum neque futurum est 16:17 et tulisti uasa decoris tui de auro meo et argento meo 16:17 Thou hast also taken thy fair jewels of my gold and

of my silver, which I had given thee, and madest to thyself images of men, and didst commit whoredom with them,

quae dedi tibi et fecisti tibi imagines masculinas et fornicata es in eis 16:18 et sumpsisti uestimenta tua multicoloria et 16:18 And tookest thy broidered garments, and coveredst

them: and thou hast set mine oil and mine incense before them.

uestita es eis et oleum meum et thymiama meum posuisti

coram eis 16:19 et panem meum quem dedi tibi similam 16:19 My meat also which I gave thee, fine flour, and oil, and honey, wherewith I fed thee, thou hast even set it before them for a sweet savour: and thus it was, saith the

et oleum et mel quibus enutriui te posuisti in conspectu eo- Lord GOD. rum in odorem suauitatis et factum est ait Dominus Deus 16:20 et tulisti filios tuos et filias tuas quas generasti mihi et 16:20 Moreover thou hast taken thy sons and thy daugh-

ters, whom thou hast borne unto me, and these hast thou sacrificed unto them to be devoured. Is this of thy whore-

immolasti eis ad deuorandum numquid parua est fornicatio doms a small matter, tua 16:21 immolantis filios meos et dedisti illos consecrans 16:21 That thou hast slain my children, and delivered them

to cause them to pass through the fire for them?

eis 16:22 et post omnes abominationes tuas et fornicatio- 16:22 And in all thine abominations and thy whoredoms

thou hast not remembered the days of thy youth, when thou wast naked and bare, and wast polluted in thy blood.

nes non es recordata dierum adulescentiae tuae quando eras nuda et confusione plena conculcata in sanguine tuo 16:23 16:23 And it came to pass after all thy wickedness, (woe,

woe unto thee! saith the LORD GOD;)

et accidit post omnem malitiam tuam uae uae tibi ait Dominus Deus 16:24 et aedificasti tibi lupanar et fecisti tibi 16:24 That thou hast also built unto thee an eminent place,

and hast made thee an high place in every street.

prostibulum in cunctis plateis 16:25 ad omne caput uiae ae- 16:25 Thou hast built thy high place at every head of the

way, and hast made thy beauty to be abhorred, and hast opened thy feet to every one that passed by, and multiplied

dificasti signum prostitutionis tuae et abominabilem fecisti thy whoredoms. decorem tuum et diuisisti pedes tuos omni transeunti et multiplicasti fornicationes tuas 16:26 et fornicata es cum filiis 16:26 Thou hast also committed fornication with the

Egyptians thy neighbours, great of flesh; and hast increased thy whoredoms, to provoke me to anger.

Aegypti uicinis tuis magnarum carnium et multiplicasti for-

nicationem tuam ad inritandum me 16:27 ecce ego extendi 16:27 Behold, therefore I have stretched out my hand over thee, and have diminished thine ordinary food, and delivered thee unto the will of them that hate thee, the daughters

manum meam super te et auferam ius tuum et dabo te in ani- of the Philistines, which are ashamed of thy lewd way. mam odientium te filiarum Palestinarum quae erubescunt in

Biblia Sacra Vulgata

16:28 Thou hast played the whore also with the Assyrians, uia tua scelerata 16:28 et fornicata es in filiis Assyriorum

because thou wast unsatiable; yea, thou hast played the harlot with them, and yet couldest not be satisfied.

eo quod necdum fueris expleta et postquam fornicata es nec

16:29 Thou hast moreover multiplied thy fornication in sic es satiata 16:29 et multiplicasti fornicationem tuam in

the land of Canaan unto Chaldea; and yet thou wast not satisfied therewith.

terra Chanaan cum Chaldeis et nec sic satiata es 16:30 in

16:30 How weak is thine heart, saith the LORD GOD, seeing thou doest all these things, the work of an imperious whorish woman;

quo mundabo cor tuum ait Dominus Deus cum facias omnia

16:31 In that thou buildest thine eminent place in the head haec opera mulieris meretricis et procacis 16:31 quia fab-

of every way, and makest thine high place in every street; and hast not been as an harlot, in that thou scornest hire;

ricasti lupanar tuum in capite omnis uiae et excelsum tuum fecisti in omni platea nec facta es quasi meretrix fastidio au-

16:32 But as a wife that committeth adultery, which taketh gens pretium 16:32 sed quasi mulier adultera quae super ui-

strangers instead of her husband!

16:33 They give gifts to all whores: but thou givest thy rum suum inducit alienos 16:33 omnibus meretricibus dan-

gifts to all thy lovers, and hirest them, that they may come unto thee on every side for thy whoredom.

tur mercedes tu autem dedisti mercedes cunctis amatoribus tuis et donabas eis ut intrarent ad te undique ad fornicandum

16:34 And the contrary is in thee from other women in tecum 16:34 factumque in te est contra consuetudinem mu-

thy whoredoms, whereas none followeth thee to commit whoredoms: and in that thou givest a reward, and no re-

ward is given unto thee, therefore thou art contrary. lierum in fornicationibus tuis et post te non erit fornicatio

in eo enim quod dedisti mercedes et mercedes non accepisti

16:35 Wherefore, O harlot, hear the word of the LORD: factum est in te contrarium 16:35 propterea meretrix audi 16:36 Thus saith the Lord GOD; Because thy filthiness uerbum Domini 16:36 haec dicit Dominus Deus quia effu-

was poured out, and thy nakedness discovered through thy whoredoms with thy lovers, and with all the idols of thy abominations, and by the blood of thy children, which thou didst give unto them;

sum est aes tuum et reuelata est ignominia tua in fornicationibus tuis super amatores tuos et super idola abominationum

16:37 Behold, therefore I will gather all thy lovers, with tuarum in sanguine filiorum tuorum quos dedisti eis 16:37 whom thou hast taken pleasure, and all them that thou hast loved, with all them that thou hast hated; I will even gather them round about against thee, and will discover thy nakedness unto them, that they may see all thy nakedness.

ecce ego congregabo omnes amatores tuos quibus commixta es et omnes quos dilexisti cum uniuersis quos oderas et congregabo eos super te undique et nudabo ignominiam tuam

16:38 And I will judge thee, as women that break wedlock coram eis et uidebunt omnem turpitudinem tuam 16:38 et and shed blood are judged; and I will give thee blood in fury and jealousy.

iudicabo te iudiciis adulterarum et effundentium sanguinem

16:39 And I will also give thee into their hand, and they et dabo te in sanguinem furoris et zeli 16:39 et dabo te in

shall throw down thine eminent place, and shall break down thy high places: they shall strip thee also of thy clothes, and shall take thy fair jewels, and leave thee naked and bare.

manus eorum et destruent lupanar tuum et demolientur prostibulum tuum et denudabunt te uestimentis tuis et auferent uasa decoris tui et derelinquent te nudam plenamque igno-

16:40 They shall also bring up a company against thee, and minia 16:40 et adducent super te multitudinem et lapidathey shall stone thee with stones, and thrust thee through with their swords.

bunt te lapidibus et trucidabunt te gladiis suis 16:41 et con-

16:41 And they shall burn thine houses with fire, and execute judgments upon thee in the sight of many women: and I will cause thee to cease from playing the harlot, and thou also shalt give no hire any more.

burent domos tuas igni et facient in te iudicia in oculis mulierum plurimarum et desines fornicari et mercedes ultra non

Prophetia Ezechielis

dabis 16:42 et requiescet indignatio mea in te et auferetur 16:42 So will I make my fury toward thee to rest, and my

jealousy shall depart from thee, and I will be quiet, and will be no more angry.

zelus meus a te et quiescam nec irascar amplius 16:43 eo

16:43 Because thou hast not remembered the days of thy youth, but hast fretted me in all these things; behold, therefore I also will recompense thy way upon thine head, saith the Lord GOD: and thou shalt not commit this lewdness above all thine abominations.

quod non fueris recordata dierum adulescentiae tuae et prouocasti me in omnibus his quapropter et ego uias tuas in capite tuo dedi ait Dominus Deus et non feci iuxta scelera tua in omnibus abominationibus tuis 16:44 ecce omnis qui dicit 16:44 Behold, every one that useth proverbs shall use this

proverb against thee, saying, As is the mother, so is her daughter.

uulgo prouerbium in te adsumet illud dicens sicut mater ita et filia eius 16:45 filia matris tuae es tu quae proiecit uirum 16:45 Thou art thy mother’s daughter, that lotheth her hus-

band and her children; and thou art the sister of thy sisters, which lothed their husbands and their children: your

suum et filios suos et soror sororum tuarum tu quae proie- mother was an Hittite, and your father an Amorite. cerunt uiros suos et filios suos mater uestra Cetthea et pater

uester Amorreus 16:46 et soror tua maior Samaria ipsa et 16:46 And thine elder sister is Samaria, she and her daughters that dwell at thy left hand: and thy younger sister, that dwelleth at thy right hand, is Sodom and her daughters.

filiae eius quae habitat ad sinistram tuam soror autem tua minor te quae habitat a dextris tuis Sodoma et filiae eius 16:47 16:47 Yet hast thou not walked after their ways, nor done

after their abominations: but, as if that were a very little thing, thou wast corrupted more than they in all thy ways.

sed nec in uiis earum ambulasti neque secundum scelera earum fecisti pauxillum minus paene sceleratiora fecisti illis in omnibus uiis tuis 16:48 uiuo ego dicit Dominus Deus quia 16:48 As I live, saith the Lord GOD, Sodom thy sister

hath not done, she nor her daughters, as thou hast done, thou and thy daughters.

non fecit Sodoma soror tua ipsa et filiae eius sicut fecisti tu et filiae tuae 16:49 ecce haec fuit iniquitas Sodomae soro- 16:49 Behold, this was the iniquity of thy sister Sodom,

pride, fulness of bread, and abundance of idleness was in her and in her daughters, neither did she strengthen the

ris tuae superbia saturitas panis et abundantia et otium ipsius hand of the poor and needy. et filiarum eius et manum egeno et pauperi non porrigebant 16:50 et eleuatae sunt et fecerunt abominationes coram me 16:50 And they were haughty, and committed abomination

before me: therefore I took them away as I saw good.

et abstuli eas sicut uidisti 16:51 et Samaria dimidium pec- 16:51 Neither hath Samaria committed half of thy sins;

but thou hast multiplied thine abominations more than they, and hast justified thy sisters in all thine abominations

catorum tuorum non peccauit sed uicisti eas sceleribus tuis which thou hast done. et iustificasti sorores tuas in omnibus abominationibus tuis

quas operata es 16:52 ergo et tu porta confusionem tuam 16:52 Thou also, which hast judged thy sisters, bear thine own shame for thy sins that thou hast committed more abominable than they: they are more righteous than thou: yea, be thou confounded also, and bear thy shame, in that thou hast justified thy sisters.

quae uicisti sorores tuas peccatis tuis sceleratius agens ab eis iustificatae sunt enim a te ergo et tu confundere et porta ig-

nominiam tuam quae iustificasti sorores tuas 16:53 et con- 16:53 When I shall bring again their captivity, the captivity of Sodom and her daughters, and the captivity of Samaria and her daughters, then will I bring again the captivity of

uertam restituens eas conuersione Sodomorum cum filiabus thy captives in the midst of them: suis et conuersione Samariae et filiarum eius et conuertam reuersionem tuam in medio earum 16:54 ut portes ignomi- 16:54 That thou mayest bear thine own shame, and mayest

be confounded in all that thou hast done, in that thou art a comfort unto them.

niam tuam et confundaris in omnibus quae fecisti consolans eas 16:55 et soror tua Sodoma et filiae eius reuertentur ad 16:55 When thy sisters, Sodom and her daughters, shall

return to their former estate, and Samaria and her daughters shall return to their former estate, then thou and thy daughters shall return to your former estate.

Biblia Sacra Vulgata

antiquitatem suam et Samaria et filiae eius reuertentur ad antiquitatem suam et tu et filiae tuae reuertimini ad antiquita-

16:56 For thy sister Sodom was not mentioned by thy tem uestram 16:56 non fuit autem Sodoma soror tua audita

mouth in the day of thy pride,

16:57 Before thy wickedness was discovered, as at the in ore tuo in die superbiae tuae 16:57 antequam reuelaretur

time of thy reproach of the daughters of Syria, and all that are round about her, the daughters of the Philistines, which

despise thee round about. malitia tua sicut hoc tempore in obprobrium filiarum Syriae

et cunctarum in circuitu tuo filiarum Palestinarum quae am-

16:58 Thou hast borne thy lewdness and thine abomina- biunt te per gyrum 16:58 scelus tuum et ignominiam tuam

tions, saith the LORD.

16:59 For thus saith the Lord GOD; I will even deal with tu portasti ait Dominus Deus 16:59 quia haec dicit Domi-

thee as thou hast done, which hast despised the oath in breaking the covenant.

nus Deus et faciam tibi sicut dispexisti iuramentum ut irri-

16:60 Nevertheless I will remember my covenant with tum faceres pactum 16:60 et recordabor ego pacti mei te-

thee in the days of thy youth, and I will establish unto thee an everlasting covenant.

cum in diebus adulescentiae tuae et suscitabo tibi pactum

16:61 Then thou shalt remember thy ways, and be sempiternum 16:61 et recordaberis uiarum tuarum et con-

ashamed, when thou shalt receive thy sisters, thine elder and thy younger: and I will give them unto thee for daugh-

ters, but not by thy covenant. funderis cum receperis sorores tuas te maiores cum minori-

16:62 And I will establish my covenant with thee; and thou bus tuis et dabo eas tibi in filias sed non ex pacto tuo 16:62 shalt know that I am the LORD:

et suscitabo ego pactum meum tecum et scies quia ego Do-

16:63 That thou mayest remember, and be confounded, minus 16:63 ut recorderis et confundaris et non sit tibi ultra

and never open thy mouth any more because of thy shame, when I am pacified toward thee for all that thou hast done,

saith the Lord GOD. aperire os prae confusione tua cum placatus fuero tibi in omnibus quae fecisti ait Dominus Deus

17:1 And the word of the LORD came unto me, saying, 17:1 et factum est uerbum Domini ad me dicens 17:2 fili

17:2 Son of man, put forth a riddle, and speak a parable unto the house of Israel;

hominis propone enigma et narra parabolam ad domum Is-

17:3 And say, Thus saith the Lord GOD; A great eagle rahel 17:3 et dices haec dicit Dominus Deus aquila gran-

with great wings, longwinged, full of feathers, which had divers colours, came unto Lebanon, and took the highest

branch of the cedar: dis magnarum alarum longo membrorum ductu plena plu-

mis et uarietate uenit ad Libanum et tulit medullam cedri

17:4 He cropped off the top of his young twigs, and carried 17:4 summitatem frondium eius auellit et transportauit eam

it into a land of traffick; he set it in a city of merchants.

17:5 He took also of the seed of the land, and planted it in in terram Chanaan in urbem negotiatorum posuit illam 17:5

a fruitful field; he placed it by great waters, and set it as a willow tree.

et tulit de semente terrae et posuit illud in terra pro semine ut firmaret radicem super aquas multas in superficie posuit il-

17:6 And it grew, and became a spreading vine of low lud 17:6 cumque germinasset creuit in uineam latiorem hu-

stature, whose branches turned toward him, and the roots thereof were under him: so it became a vine, and brought

forth branches, and shot forth sprigs. mili statura respicientibus ramis eius ad eam et radices eius

sub illa erunt facta est ergo uinea et fructificauit in palmites

17:7 There was also another great eagle with great wings et emisit propagines 17:7 et facta est aquila altera grandis

and many feathers: and, behold, this vine did bend her roots toward him, and shot forth her branches toward him,

that he might water it by the furrows of her plantation. magnis alis multisque plumis et ecce uinea ista quasi mittens radices suas ad eam palmites suos extendit ad illam ut inri-

Prophetia Ezechielis

garet eam de areolis germinis sui 17:8 in terra bona super 17:8 It was planted in a good soil by great waters, that it

might bring forth branches, and that it might bear fruit, that it might be a goodly vine.

aquas multas plantata est ut faciat frondes et portet fructum et sit in uineam grandem 17:9 dic haec dicit Dominus Deus 17:9 Say thou, Thus saith the Lord GOD; Shall it prosper?

shall he not pull up the roots thereof, and cut off the fruit thereof, that it wither? it shall wither in all the leaves of her spring, even without great power or many people to pluck it up by the roots thereof.

ergone prosperabitur nonne radices eius euellet et fructum eius distringet et siccabit omnes palmites germinis eius et arescet et non in brachio grandi neque in populo multo ut euelleret eam radicitus 17:10 ecce plantata est ergone pros- 17:10 Yea, behold, being planted, shall it prosper? shall it

not utterly wither, when the east wind toucheth it? it shall wither in the furrows where it grew.

perabitur nonne cum tetigerit eam uentus urens siccabitur et in areis germinis sui arescet 17:11 et factum est uerbum Do- 17:11 Moreover the word of the LORD came unto me,

saying,

mini ad me dicens 17:12 dic ad domum exasperantem nes- 17:12 Say now to the rebellious house, Know ye not what

these things mean? tell them, Behold, the king of Babylon is come to Jerusalem, and hath taken the king thereof, and

citis quid ista significent dic ecce uenit rex Babylonis Hieru- the princes thereof, and led them with him to Babylon; salem et adsumet regem et principes eius et adducet eos ad semet ipsum in Babylonem 17:13 et tollet de semine regni 17:13 And hath taken of the king’s seed, and made a

covenant with him, and hath taken an oath of him: he hath also taken the mighty of the land:

ferietque cum eo foedus et accipiet ab eo iusiurandum sed et fortes terrae tollet 17:14 ut sit regnum humile et non eleue- 17:14 That the kingdom might be base, that it might not

lift itself up, but that by keeping of his covenant it might stand.

tur sed custodiat pactum eius et seruet illud 17:15 qui rece-

17:15 But he rebelled against him in sending his ambassadors into Egypt, that they might give him horses and much people. Shall he prosper? shall he escape that doeth such things? or shall he break the covenant, and be delivered?

dens ab eo misit nuntios ad Aegyptum ut daret sibi equos et populum multum numquid prosperabitur uel consequetur salutem qui fecit haec et qui dissoluit pactum numquid effugiet 17:16 uiuo ego dicit Dominus Deus quoniam in loco regis qui 17:16 As I live, saith the Lord GOD, surely in the place

where the king dwelleth that made him king, whose oath he despised, and whose covenant he brake, even with him

constituit eum regem cuius fecit irritum iuramentum et soluit in the midst of Babylon he shall die. pactum quod habebat cum eo in medio Babylonis morietur 17:17 et non in exercitu grandi neque in populo multo faciet 17:17 Neither shall Pharaoh with his mighty army and

great company make for him in the war, by casting up mounts, and building forts, to cut off many persons:

contra eum Pharao proelium in iactu aggeris et in extructione uallorum ut interficiat animas multas 17:18 spreuerat enim 17:18 Seeing he despised the oath by breaking the

covenant, when, lo, he had given his hand, and hath done all these things, he shall not escape.

iuramentum ut solueret foedus et ecce dedit manum suam et cum omnia haec fecerit non effugiet 17:19 propterea haec 17:19 Therefore thus saith the Lord GOD; As I live, surely

mine oath that he hath despised, and my covenant that he hath broken, even it will I recompense upon his own head.

dicit Dominus Deus uiuo ego quoniam iuramentum quod spreuit et foedus quod praeuaricatus est ponam in caput eius 17:20 et expandam super eum rete meum et conprehendetur 17:20 And I will spread my net upon him, and he shall

be taken in my snare, and I will bring him to Babylon, and will plead with him there for his trespass that he hath

sagena mea et adducam eum in Babylonem et iudicabo illum trespassed against me. ibi in praeuaricatione qua despexit me 17:21 et omnes pro- 17:21 And all his fugitives with all his bands shall fall by

the sword, and they that remain shall be scattered toward all winds: and ye shall know that I the LORD have spoken

fugi eius cum uniuerso agmine gladio cadent residui autem it.

Biblia Sacra Vulgata

in omnem uentum dispergentur et scietis quia ego Dominus

17:22 Thus saith the Lord GOD; I will also take of the locutus sum 17:22 haec dicit Dominus Deus et sumam ego

highest branch of the high cedar, and will set it; I will crop off from the top of his young twigs a tender one, and will

plant it upon an high mountain and eminent: de medulla cedri sublimis et ponam de uertice ramorum eius

tenerum distringam et plantabo super montem excelsum et

17:23 In the mountain of the height of Israel will I plant eminentem 17:23 in monte sublimi Israhel plantabo illud et

it: and it shall bring forth boughs, and bear fruit, and be a goodly cedar: and under it shall dwell all fowl of every wing; in the shadow of the branches thereof shall they dwell.

erumpet in germen et faciet fructum et erit in cedrum magnam et habitabunt sub eo omnes uolucres uniuersum uolatile

17:24 And all the trees of the field shall know that I the sub umbra frondium eius nidificabit 17:24 et scient omnia

LORD have brought down the high tree, have exalted the low tree, have dried up the green tree, and have made the dry tree to flourish: I the LORD have spoken and have done it.

ligna regionis quia ego Dominus humiliaui lignum sublime et exaltaui lignum humile et siccaui lignum uiride et frondere feci lignum aridum ego Dominus locutus sum et feci

18:1 The word of the LORD came unto me again, saying, 18:1 et factus est sermo Domini ad me dicens 18:2 quid

18:2 What mean ye, that ye use this proverb concerning the land of Israel, saying, The fathers have eaten sour grapes, and the children’s teeth are set on edge?

est quod inter uos parabolam uertitis in prouerbium istud in terra Israhel dicentes patres comederunt uuam acerbam et

18:3 As I live, saith the Lord GOD, ye shall not have oc- dentes filiorum obstupescunt 18:3 uiuo ego dicit Dominus

casion any more to use this proverb in Israel.

Deus si erit uobis ultra parabola haec in prouerbium in Is-

18:4 Behold, all souls are mine; as the soul of the father, rahel 18:4 ecce omnes animae meae sunt ut anima patris so also the soul of the son is mine: the soul that sinneth, it shall die.

ita et anima filii mea est anima quae peccauerit ipsa morie-

18:5 But if a man be just, and do that which is lawful and tur 18:5 et uir si fuerit iustus et fecerit iudicium et iustitiam

right,

18:6 And hath not eaten upon the mountains, neither hath 18:6 in montibus non comederit et oculos suos non leuauerit

lifted up his eyes to the idols of the house of Israel, neither hath defiled his neighbour’s wife, neither hath come near

to a menstruous woman, ad idola domus Israhel et uxorem proximi sui non uiolauerit

18:7 And hath not oppressed any, but hath restored to the et ad mulierem menstruatam non accesserit 18:7 et homidebtor his pledge, hath spoiled none by violence, hath given his bread to the hungry, and hath covered the naked

with a garment; nem non contristauerit pignus debitori reddiderit per uim ni-

hil rapuerit panem suum esurienti dederit et nudum operue-

18:8 He that hath not given forth upon usury, neither hath rit uestimento 18:8 ad usuram non commodauerit et amptaken any increase, that hath withdrawn his hand from iniquity, hath executed true judgment between man and man,

lius non acceperit ab iniquitate auerterit manum suam iudi-

18:9 Hath walked in my statutes, and hath kept my judg- cium uerum fecerit inter uirum et uirum 18:9 in praeceptis

ments, to deal truly; he is just, he shall surely live, saith the Lord GOD.

meis ambulauerit et iudicia mea custodierit ut faciat uerita-

18:10 If he beget a son that is a robber, a shedder of blood, tem hic iustus est uita uiuet ait Dominus Deus 18:10 quod and that doeth the like to any one of these things,

si genuerit filium latronem effundentem sanguinem et fece-

18:11 And that doeth not any of those duties, but even rit unum de istis 18:11 et haec quidem omnia non facientem

hath eaten upon the mountains, and defiled his neighbour’s wife,

sed in montibus comedentem et uxorem proximi sui polluen-

18:12 Hath oppressed the poor and needy, hath spoiled by tem 18:12 egenum et pauperem contristantem rapientem ra-

violence, hath not restored the pledge, and hath lifted up his eyes to the idols, hath committed abomination,

Prophetia Ezechielis

pinas pignus non reddentem et ad idola leuantem oculos suos abominationem facientem 18:13 ad usuram dantem et amp- 18:13 Hath given forth upon usury, and hath taken in-

crease: shall he then live? he shall not live: he hath done all these abominations; he shall surely die; his blood shall

lius accipientem numquid uiuet non uiuet cum uniuersa de- be upon him. testanda haec fecerit morte morietur sanguis eius in ipso erit 18:14 quod si genuerit filium qui uidens omnia peccata patris 18:14 Now, lo, if he beget a son, that seeth all his father’s

sins which he hath done, and considereth, and doeth not such like,

sui quae fecit timuerit et non fecerit simile eis 18:15 super

18:15 That hath not eaten upon the mountains, neither hath lifted up his eyes to the idols of the house of Israel, hath not defiled his neighbour’s wife,

montes non comederit et oculos suos non leuauerit ad idola

domus Israhel et uxorem proximi sui non uiolauerit 18:16 18:16 Neither hath oppressed any, hath not withholden the pledge, neither hath spoiled by violence, but hath given his bread to the hungry, and hath covered the naked with a

et uirum non contristauerit pignus non retinuerit et rapinam garment, non rapuerit panem suum esurienti dederit et nudum operuerit uestimento 18:17 a pauperis iniuria auerterit manum 18:17 That hath taken off his hand from the poor, that hath

not received usury nor increase, hath executed my judgments, hath walked in my statutes; he shall not die for the

suam usuram et superabundantiam non acceperit iudicia mea iniquity of his father, he shall surely live. fecerit in praeceptis meis ambulauerit hic non morietur in iniquitate patris sui sed uita uiuet 18:18 pater eius quia calum- 18:18 As for his father, because he cruelly oppressed,

spoiled his brother by violence, and did that which is not good among his people, lo, even he shall die in his iniq-

niatus est et uim fecit fratri et malum operatus est in medio uity. populi sui ecce mortuus est in iniquitate sua 18:19 et dicitis 18:19 Yet say ye, Why? doth not the son bear the iniquity

of the father? When the son hath done that which is lawful and right, and hath kept all my statutes, and hath done

quare non portauit filius iniquitatem patris uidelicet quia fi- them, he shall surely live. lius iudicium et iustitiam operatus est omnia praecepta mea custodiuit et fecit illa uita uiuet 18:20 anima quae pecca- 18:20 The soul that sinneth, it shall die. The son shall not

bear the iniquity of the father, neither shall the father bear the iniquity of the son: the righteousness of the righteous shall be upon him, and the wickedness of the wicked shall be upon him.

uerit ipsa morietur filius non portabit iniquitatem patris et pater non portabit iniquitatem filii iustitia iusti super eum erit et impietas impii erit super eum 18:21 si autem impius 18:21 But if the wicked will turn from all his sins that

he hath committed, and keep all my statutes, and do that which is lawful and right, he shall surely live, he shall not

egerit paenitentiam ab omnibus peccatis suis quae operatus die. est et custodierit uniuersa praecepta mea et fecerit iudicium et iustitiam uita uiuet non morietur 18:22 omnium iniqui- 18:22 All his transgressions that he hath committed, they

shall not be mentioned unto him: in his righteousness that he hath done he shall live.

tatum eius quas operatus est non recordabor in iustitia sua quam operatus est uiuet 18:23 numquid uoluntatis meae est 18:23 Have I any pleasure at all that the wicked should

die? saith the Lord GOD: and not that he should return from his ways, and live?

mors impii dicit Dominus Deus et non ut conuertatur a uiis suis et uiuat 18:24 si autem auerterit se iustus a iustitia sua 18:24 But when the righteous turneth away from his righ-

teousness, and committeth iniquity, and doeth according to all the abominations that the wicked man doeth, shall he live? All his righteousness that he hath done shall not be mentioned: in his trespass that he hath trespassed, and in his sin that he hath sinned, in them shall he die.

et fecerit iniquitatem secundum omnes abominationes quas operari solet impius numquid uiuet omnes iustitiae eius quas fecerat non recordabuntur in praeuaricatione qua praeuaricatus est et in peccato suo quod peccauit in ipsis morietur

Biblia Sacra Vulgata

18:25 Yet ye say, The way of the LORD is not equal. Hear 18:25 et dixistis non est aequa uia Domini audite domus Is-

now, O house of Israel; Is not my way equal? are not your ways unequal?

rahel numquid uia mea non est aequa et non magis uiae ue-

18:26 When a righteous man turneth away from his righ- strae prauae sunt 18:26 cum enim auerterit se iustus a iusti-

teousness, and committeth iniquity, and dieth in them; for his iniquity that he hath done shall he die.

tia sua et fecerit iniquitatem morietur in eis in iniustitia quam

18:27 Again, when the wicked man turneth away from his operatus est morietur 18:27 et cum auerterit se impius ab

wickedness that he hath committed, and doeth that which is lawful and right, he shall save his soul alive.

impietate sua quam operatus est et fecerit iudicium et iusti-

18:28 Because he considereth, and turneth away from all tiam ipse animam suam uiuificabit 18:28 considerans enim

his transgressions that he hath committed, he shall surely live, he shall not die.

et auertens se ab omnibus iniquitatibus suis quas operatus

18:29 Yet saith the house of Israel, The way of the LORD est uita uiuet et non morietur 18:29 et dicunt filii Israhel

is not equal. O house of Israel, are not my ways equal? are not your ways unequal?

non est aequa uia Domini numquid uiae meae non sunt ae-

18:30 Therefore I will judge you, O house of Israel, every quae domus Israhel et non magis uiae uestrae prauae 18:30 one according to his ways, saith the Lord GOD. Repent, and turn yourselves from all your transgressions; so iniq-

uity shall not be your ruin. idcirco unumquemque iuxta uias suas iudicabo domus Israhel ait Dominus Deus conuertimini et agite paenitentiam ab omnibus iniquitatibus uestris et non erit uobis in ruinam

18:31 Cast away from you all your transgressions, iniquitas 18:31 proicite a uobis omnes praeuaricationes ue-

whereby ye have transgressed; and make you a new heart and a new spirit: for why will ye die, O house of Israel?

stras in quibus praeuaricati estis et facite uobis cor nouum

18:32 For I have no pleasure in the death of him that dieth, et spiritum nouum et quare moriemini domus Israhel 18:32 saith the Lord GOD: wherefore turn yourselves, and live ye.

quia nolo mortem morientis dicit Dominus Deus reuertimini et uiuite

19:1 Moreover take thou up a lamentation for the princes 19:1 et tu adsume planctum super principes Israhel 19:2 of Israel, 19:2 And say, What is thy mother? A lioness: she lay down among lions, she nourished her whelps among young lions.

et dices quare mater tua leaena inter leones cubauit in me-

19:3 And she brought up one of her whelps: it became a dio leunculorum enutriuit catulos suos 19:3 et eduxit unum

young lion, and it learned to catch the prey; it devoured men.

de leunculis suis leo factus est et didicit capere praedam ho-

19:4 The nations also heard of him; he was taken in their minemque comedere 19:4 et audierunt de eo gentes et non pit, and they brought him with chains unto the land of Egypt.

absque uulneribus suis ceperunt eum et adduxerunt eum in

19:5 Now when she saw that she had waited, and her hope catenis in terram Aegypti 19:5 quae cum uidisset quoniam

was lost, then she took another of her whelps, and made him a young lion.

infirmata est et periit expectatio eius tulit unum de leunculis

19:6 And he went up and down among the lions, he be- suis leonem constituit eum 19:6 qui incedebat inter leones

came a young lion, and learned to catch the prey, and devoured men.

et factus est leo didicit praedam capere et homines deuo-

19:7 And he knew their desolate palaces, and he laid waste rare 19:7 didicit uiduas facere et ciuitates eorum in deser-

their cities; and the land was desolate, and the fulness thereof, by the noise of his roaring.

tum adducere et desolata est terra et plenitudo eius a uoce

19:8 Then the nations set against him on every side from rugitus illius 19:8 et conuenerunt aduersum eum gentes un-

the provinces, and spread their net over him: he was taken in their pit.

dique de prouinciis et expanderunt super eum rete suum in

Prophetia Ezechielis

uulneribus earum captus est 19:9 et miserunt eum in ca- 19:9 And they put him in ward in chains, and brought him

to the king of Babylon: they brought him into holds, that his voice should no more be heard upon the mountains of

ueam in catenis adduxerunt eum ad regem Babylonis mi- Israel. seruntque eum in carcerem ne audiretur uox eius ultra super montes Israhel 19:10 mater tua quasi uinea in sanguine 19:10 Thy mother is like a vine in thy blood, planted by

the waters: she was fruitful and full of branches by reason of many waters.

tuo super aquam plantata fructus eius et frondes eius creuerunt ex aquis multis 19:11 et factae sunt ei uirgae soli- 19:11 And she had strong rods for the sceptres of them

that bare rule, and her stature was exalted among the thick branches, and she appeared in her height with the multi-

dae in sceptra dominantium et exaltata est statura eius inter tude of her branches. frondes et uidit altitudinem suam in multitudine palmitum suorum 19:12 et euulsa est in ira in terramque proiecta et 19:12 But she was plucked up in fury, she was cast down

to the ground, and the east wind dried up her fruit: her strong rods were broken and withered; the fire consumed

uentus urens siccauit fructum eius marcuerunt et arefactae them. sunt uirgae roboris eius ignis comedit eam 19:13 et nunc 19:13 And now she is planted in the wilderness, in a dry

and thirsty ground.

transplantata est in desertum in terra inuia et sitienti 19:14 19:14 And fire is gone out of a rod of her branches, which

hath devoured her fruit, so that she hath no strong rod to be a sceptre to rule. This is a lamentation, and shall be for

et egressus est ignis de uirga ramorum eius qui fructum eius a lamentation. comedit et non fuit in ea uirga fortis sceptrum dominantium planctus est et erit in planctum

20:1 et factum est in anno septimo in quinto mense in de- 20:1 And it came to pass in the seventh year, in the fifth

month, the tenth day of the month, that certain of the elders of Israel came to enquire of the LORD, and sat before me.

cima mensis uenerunt uiri de senioribus Israhel ut interrogarent Dominum et sederunt coram me 20:2 et factus est 20:2 Then came the word of the LORD unto me, saying,

sermo Domini ad me dicens 20:3 fili hominis loquere se- 20:3 Son of man, speak unto the elders of Israel, and say unto them, Thus saith the Lord GOD; Are ye come to enquire of me? As I live, saith the Lord GOD, I will not be

nioribus Israhel et dices ad eos haec dicit Dominus Deus enquired of by you. num ad interrogandum me uos uenistis uiuo ego quia non respondebo uobis ait Dominus Deus 20:4 si iudicas eos si 20:4 Wilt thou judge them, son of man, wilt thou judge

them? cause them to know the abominations of their fathers:

iudicas fili hominis abominationes patrum eorum ostende eis 20:5 et dices ad eos haec dicit Dominus Deus in die qua elegi 20:5 And say unto them, Thus saith the Lord GOD; In the

day when I chose Israel, and lifted up mine hand unto the seed of the house of Jacob, and made myself known unto them in the land of Egypt, when I lifted up mine hand unto them, saying, I am the LORD your God;

Israhel et leuaui manum meam pro stirpe domus Iacob et apparui eis in terra Aegypti et leuaui manum meam pro eis dicens ego Dominus Deus uester 20:6 in die illa leuaui ma- 20:6 In the day that I lifted up mine hand unto them, to

bring them forth of the land of Egypt into a land that I had espied for them, flowing with milk and honey, which is the

num meam pro eis ut educerem eos de terra Aegypti in ter- glory of all lands: ram quam prouideram eis fluentem lacte et melle quae est egregia inter omnes terras 20:7 et dixi ad eos unusquisque 20:7 Then said I unto them, Cast ye away every man the

abominations of his eyes, and defile not yourselves with the idols of Egypt: I am the LORD your God.

offensiones oculorum suorum abiciat et in idolis Aegypti nolite pollui ego Dominus Deus uester 20:8 et inritauerunt me 20:8 But they rebelled against me, and would not hearken

unto me: they did not every man cast away the abominations of their eyes, neither did they forsake the idols of Egypt: then I said, I will pour out my fury upon them, to accomplish my anger against them in the midst of the land of Egypt.

nolueruntque audire unusquisque abominationes oculorum

Biblia Sacra Vulgata

suorum non proiecit nec idola Aegypti reliquerunt et dixi ut effunderem indignationem meam super eos et implerem

20:9 But I wrought for my name’s sake, that it should not iram meam in eis in medio terrae Aegypti 20:9 et feci pro-

be polluted before the heathen, among whom they were, in whose sight I made myself known unto them, in bringing

them forth out of the land of Egypt. pter nomen meum ut non uiolaretur coram gentibus in qua-

rum medio erant et inter quas apparui eis ut educerem eos

20:10 Wherefore I caused them to go forth out of the land de terra Aegypti 20:10 eieci ergo eos de terra Aegypti et

of Egypt, and brought them into the wilderness.

20:11 And I gave them my statutes, and shewed them my eduxi in desertum 20:11 et dedi eis praecepta mea et iudi-

judgments, which if a man do, he shall even live in them.

20:12 Moreover also I gave them my sabbaths, to be a sign cia mea ostendi eis quae faciat homo et uiuat in eis 20:12

between me and them, that they might know that I am the LORD that sanctify them.

insuper et sabbata mea dedi eis ut esset signum inter me et

20:13 But the house of Israel rebelled against me in the eos et scirent quia ego Dominus sanctificans eos 20:13 et

wilderness: they walked not in my statutes, and they despised my judgments, which if a man do, he shall even live in them; and my sabbaths they greatly polluted: then I said, I would pour out my fury upon them in the wilderness, to consume them.

inritauerunt me domus Israhel in deserto in praeceptis meis non ambulauerunt et iudicia mea proiecerunt quae faciens homo uiuet in eis et sabbata mea uiolauerunt uehementer dixi ergo ut effunderem furorem meum super eos in deserto

20:14 But I wrought for my name’s sake, that it should not et consumerem eos 20:14 et feci propter nomen meum ne be polluted before the heathen, in whose sight I brought them out.

uiolaretur coram gentibus de quibus eieci eos in conspectu

20:15 Yet also I lifted up my hand unto them in the wilder- earum 20:15 ego igitur leuaui manum meam super eos in ness, that I would not bring them into the land which I had given them, flowing with milk and honey, which is the

glory of all lands; deserto ne inducerem eos in terram quam dedi eis fluentem 20:16 Because they despised my judgments, and walked lacte et melle praecipuam terrarum omnium 20:16 quia iu-

not in my statutes, but polluted my sabbaths: for their heart went after their idols.

dicia mea proiecerunt et in praeceptis meis non ambulauerunt et sabbata mea uiolauerunt post idola enim cor eorum

20:17 Nevertheless mine eye spared them from destroying gradiebatur 20:17 et pepercit oculus meus super eos ut non

them, neither did I make an end of them in the wilderness. 20:18 But I said unto their children in the wilderness, Walk interficerem eos nec consumpsi eos in deserto 20:18 dixi ye not in the statutes of your fathers, neither observe their judgments, nor defile yourselves with their idols:

autem ad filios eorum in solitudine in praeceptis patrum uestrorum nolite incedere nec iudicia eorum custodiatis nec in

20:19 I am the LORD your God; walk in my statutes, and idolis eorum polluamini 20:19 ego Dominus Deus uester in

keep my judgments, and do them;

praeceptis meis ambulate et iudicia mea custodite et facite

20:20 And hallow my sabbaths; and they shall be a sign ea 20:20 et sabbata mea sanctificate ut sit signum inter me

between me and you, that ye may know that I am the LORD your God.

et uos et sciatur quia ego Dominus Deus uester 20:21 et ex-

20:21 Notwithstanding the children rebelled against me: they walked not in my statutes, neither kept my judgments to do them, which if a man do, he shall even live in them; they polluted my sabbaths: then I said, I would pour out my fury upon them, to accomplish my anger against them in the wilderness.

acerbauerunt me filii in praeceptis meis non ambulauerunt et iudicia mea non custodierunt ut facerent ea quae cum fecerit homo uiuet in eis et sabbata mea uiolauerunt et comminatus sum ut effunderem furorem meum super eos et imple-

Prophetia Ezechielis

rem iram meam in eis in deserto 20:22 auerti autem ma- 20:22 Nevertheless I withdrew mine hand, and wrought for my name’s sake, that it should not be polluted in the sight of the heathen, in whose sight I brought them forth.

num meam et feci propter nomen meum ut non uiolaretur coram gentibus de quibus eieci eos in oculis earum 20:23 20:23 I lifted up mine hand unto them also in the wilder-

ness, that I would scatter them among the heathen, and disperse them through the countries;

iterum leuaui manum meam in eos in solitudine ut dispergerem illos in nationes et uentilarem in terras 20:24 eo quod 20:24 Because they had not executed my judgments, but

had despised my statutes, and had polluted my sabbaths, and their eyes were after their fathers’ idols.

iudicia mea non fecissent et praecepta mea reprobassent et sabbata mea uiolassent et post idola patrum suorum fuissent oculi eorum 20:25 ergo et ego dedi eis praecepta non bona 20:25 Wherefore I gave them also statutes that were not

good, and judgments whereby they should not live;

et iudicia in quibus non uiuent 20:26 et pollui eos in mu- 20:26 And I polluted them in their own gifts, in that they

caused to pass through the fire all that openeth the womb, that I might make them desolate, to the end that they might

neribus suis cum offerrent omne quod aperit uuluam prop- know that I am the LORD. ter delicta sua et scient quia ego Dominus 20:27 quam ob 20:27 Therefore, son of man, speak unto the house of Is-

rael, and say unto them, Thus saith the Lord GOD; Yet in this your fathers have blasphemed me, in that they have

rem loquere ad domum Israhel fili hominis et dices ad eos committed a trespass against me. haec dicit Dominus Deus adhuc et in hoc blasphemauerunt me patres uestri cum spreuissent me contemnentes 20:28 et 20:28 For when I had brought them into the land, for the

which I lifted up mine hand to give it to them, then they saw every high hill, and all the thick trees, and they offered there their sacrifices, and there they presented the provocation of their offering: there also they made their sweet savour, and poured out there their drink offerings.

induxissem eos in terram super quam leuaui manum meam ut darem eis uiderunt omnem collem excelsum et omne lignum nemorosum et immolauerunt ibi uictimas suas et dederunt ibi inritationem oblationis suae et posuerunt ibi odorem suauitatis suae et libauerunt libationes suas 20:29 et dixi ad 20:29 Then I said unto them, What is the high place where-

unto ye go? And the name whereof is called Bamah unto this day.

eos quid est excelsum ad quod uos ingredimini et uocatum

est nomen eius Excelsum usque ad hanc diem 20:30 prop- 20:30 Wherefore say unto the house of Israel, Thus saith the Lord GOD; Are ye polluted after the manner of your fathers? and commit ye whoredom after their abomina-

terea dic ad domum Israhel haec dicit Dominus Deus certe tions? in uia patrum uestrorum uos polluimini et post offendicula eorum uos fornicamini 20:31 et in oblatione donorum ue- 20:31 For when ye offer your gifts, when ye make your

sons to pass through the fire, ye pollute yourselves with all your idols, even unto this day: and shall I be enquired of by you, O house of Israel? As I live, saith the Lord GOD, I will not be enquired of by you.

strorum cum transducitis filios uestros per ignem uos polluimini in omnibus idolis uestris usque hodie et ego respondebo uobis domus Israhel uiuo ego dicit Dominus Deus quia non respondebo uobis 20:32 neque cogitatio mentis uestrae 20:32 And that which cometh into your mind shall not be

at all, that ye say, We will be as the heathen, as the families of the countries, to serve wood and stone.

fiet dicentium erimus sicut gentes et sicut cognationes terrae ut colamus ligna et lapides 20:33 uiuo ego dicit Dominus 20:33 As I live, saith the Lord GOD, surely with a mighty

hand, and with a stretched out arm, and with fury poured out, will I rule over you:

Deus quoniam in manu forti et brachio extento et in furore effuso regnabo super uos 20:34 et educam uos de populis et 20:34 And I will bring you out from the people, and will

gather you out of the countries wherein ye are scattered, with a mighty hand, and with a stretched out arm, and

congregabo uos de terris in quibus dispersi estis in manu ua- with fury poured out.

Biblia Sacra Vulgata

lida et brachio extento et in furore effuso regnabo super uos

20:35 And I will bring you into the wilderness of the peo- 20:35 et adducam uos in desertum populorum et iudicabor ple, and there will I plead with you face to face. 20:36 Like as I pleaded with your fathers in the wilderness uobiscum ibi facie ad faciem 20:36 sicut iudicio contendi of the land of Egypt, so will I plead with you, saith the Lord GOD.

aduersum patres uestros in deserto terrae Aegypti sic iudi-

20:37 And I will cause you to pass under the rod, and I cabo uos dicit Dominus Deus 20:37 et subiciam uos sceptro

will bring you into the bond of the covenant:

20:38 And I will purge out from among you the rebels, and meo et inducam uos in uinculis foederis 20:38 et eligam de

them that transgress against me: I will bring them forth out of the country where they sojourn, and they shall not enter into the land of Israel: and ye shall know that I am the LORD.

uobis transgressores et impios et de terra incolatus eorum educam eos et terram Israhel non ingredientur et scietis quia

20:39 As for you, O house of Israel, thus saith the Lord ego Dominus 20:39 et uos domus Israhel haec dicit Domi-

GOD; Go ye, serve ye every one his idols, and hereafter also, if ye will not hearken unto me: but pollute ye my

holy name no more with your gifts, and with your idols. nus Deus singuli post idola uestra ambulate et seruite eis quod si et in hoc non audieritis me et nomen meum sanctum pollueritis ultra in muneribus uestris et in idolis uestris

20:40 For in mine holy mountain, in the mountain of the 20:40 in monte sancto meo in monte excelso Israhel ait Do-

height of Israel, saith the Lord GOD, there shall all the house of Israel, all of them in the land, serve me: there will I accept them, and there will I require your offerings, and the firstfruits of your oblations, with all your holy things.

minus Deus ibi seruiet mihi omnis domus Israhel omnes inquam in terra in qua placebunt mihi et ibi quaeram primitias uestras et initium decimarum uestrarum in omnibus sanctifi-

20:41 I will accept you with your sweet savour, when I cationibus uestris 20:41 in odorem suauitatis suscipiam uos

bring you out from the people, and gather you out of the countries wherein ye have been scattered; and I will be

sanctified in you before the heathen. cum eduxero uos de populis et congregauero uos de terris in

quas dispersi estis et sanctificabor in uobis in oculis natio-

20:42 And ye shall know that I am the LORD, when I shall num 20:42 et scietis quia ego Dominus cum induxero uos

bring you into the land of Israel, into the country for the which I lifted up mine hand to give it to your fathers.

ad terram Israhel in terram pro qua leuaui manum meam ut

20:43 And there shall ye remember your ways, and all darem eam patribus uestris 20:43 et recordabimini ibi uia-

your doings, wherein ye have been defiled; and ye shall lothe yourselves in your own sight for all your evils that

ye have committed. rum uestrarum et omnium scelerum uestrorum quibus pol-

luti estis in eis et displicebitis uobis in conspectu uestro in

20:44 And ye shall know that I am the LORD when I have omnibus malitiis uestris quas fecistis 20:44 et scietis quia

wrought with you for my name’s sake, not according to your wicked ways, nor according to your corrupt doings,

O ye house of Israel, saith the Lord GOD. ego Dominus cum benefecero uobis propter nomen meum

non secundum uias uestras malas neque secundum scelera

20:45 Moreover the word of the LORD came unto me, uestra pessima domus Israhel ait Dominus Deus 20:45 et

saying, 20:46 Son of man, set thy face toward the south, and drop factus est sermo Domini ad me dicens 20:46 fili hominis thy word toward the south, and prophesy against the forest of the south field;

pone faciem tuam contra uiam austri et stilla ad africum et

20:47 And say to the forest of the south, Hear the word propheta ad saltum agri meridiani 20:47 et dices saltui me-

of the LORD; Thus saith the Lord GOD; Behold, I will kindle a fire in thee, and it shall devour every green tree in thee, and every dry tree: the flaming flame shall not be quenched, and all faces from the south to the north shall be burned therein.

ridiano audi uerbum Domini haec dicit Dominus Deus ecce ego succendam in te ignem et conburam in te omne lignum

Prophetia Ezechielis

uiride et omne lignum aridum non extinguetur flamma succensionis et conburetur in ea omnis facies ab austro usque ad aquilonem 20:48 et uidebit uniuersa caro quia ego Do- 20:48 And all flesh shall see that I the LORD have kindled

it: it shall not be quenched.

minus succendi eam nec extinguetur 20:49 et dixi ha ha ha 20:49 Then said I, Ah Lord GOD! they say of me, Doth he

not speak parables?

Domine Deus ipsi dicunt de me numquid non per parabolas loquitur iste

21:1 et factus est sermo Domini ad me dicens 21:2 fili 21:1 And the word of the LORD came unto me, saying,

21:2 Son of man, set thy face toward Jerusalem, and drop thy word toward the holy places, and prophesy against the land of Israel,

hominis pone faciem tuam ad Hierusalem et stilla ad sanctuaria et propheta contra humum Israhel 21:3 et dices terrae 21:3 And say to the land of Israel, Thus saith the LORD;

Behold, I am against thee, and will draw forth my sword out of his sheath, and will cut off from thee the righteous

Israhel haec dicit Dominus Deus ecce ego ad te et eiciam gla- and the wicked. dium meum de uagina sua et occidam in te iustum et impium 21:4 pro eo autem quod occidi in te iustum et impium idcirco 21:4 Seeing then that I will cut off from thee the righteous

and the wicked, therefore shall my sword go forth out of his sheath against all flesh from the south to the north:

egredietur gladius meus de uagina sua ad omnem carnem ab austro ad aquilonem 21:5 ut sciat omnis caro quia ego 21:5 That all flesh may know that I the LORD have drawn

forth my sword out of his sheath: it shall not return any more.

Dominus eduxi gladium meum de uagina sua inreuocabilem 21:6 et tu fili hominis ingemesce in contritione lumborum et 21:6 Sigh therefore, thou son of man, with the breaking of

thy loins; and with bitterness sigh before their eyes.

in amaritudinibus ingemesce coram eis 21:7 cumque dixe- 21:7 And it shall be, when they say unto thee, Wherefore

sighest thou? that thou shalt answer, For the tidings; because it cometh: and every heart shall melt, and all hands shall be feeble, and every spirit shall faint, and all knees shall be weak as water: behold, it cometh, and shall be brought to pass, saith the Lord GOD.

rint ad te quare tu gemis dices pro auditu quia uenit et tabescet omne cor et dissoluentur uniuersae manus et infirmabitur omnis spiritus et per cuncta genua fluent aquae ecce uenit et fiet ait Dominus Deus 21:8 et factus est sermo Domini 21:8 Again the word of the LORD came unto me, saying, ad me dicens 21:9 fili hominis propheta et dices haec dicit 21:9 Son of man, prophesy, and say, Thus saith the LORD;

Say, A sword, a sword is sharpened, and also furbished:

Dominus Deus loquere gladius gladius exacutus est et limatus 21:10 ut caedat uictimas exacutus est ut splendeat lima- 21:10 It is sharpened to make a sore slaughter; it is fur-

bished that it may glitter: should we then make mirth? it contemneth the rod of my son, as every tree.

tus est qui moues sceptrum filii mei succidisti omne lignum 21:11 et dedi eum ad leuigandum ut teneatur manu iste ex- 21:11 And he hath given it to be furbished, that it may be

handled: this sword is sharpened, and it is furbished, to give it into the hand of the slayer.

acutus est gladius et iste limatus ut sit in manu interficientis 21:12 clama et ulula fili hominis quia hic factus est in populo 21:12 Cry and howl, son of man: for it shall be upon my

people, it shall be upon all the princes of Israel: terrors by reason of the sword shall be upon my people: smite

meo hic in cunctis ducibus Israhel qui fugerant gladio tra- therefore upon thy thigh. diti sunt cum populo meo idcirco plaude super femur 21:13 21:13 Because it is a trial, and what if the sword contemn

even the rod? it shall be no more, saith the Lord GOD.

quia probatus est et hoc cum sceptrum subuerterit et non erit dicit Dominus Deus 21:14 tu ergo fili hominis propheta et 21:14 Thou therefore, son of man, prophesy, and smite

thine hands together. and let the sword be doubled the third time, the sword of the slain: it is the sword of the great men that are slain, which entereth into their privy chambers.

percute manu ad manum et duplicetur gladius ac triplicetur

Biblia Sacra Vulgata

gladius interfectorum hic est gladius occisionis magnae qui

21:15 I have set the point of the sword against all their obstupescere eos facit 21:15 et corde tabescere et multipli-

gates, that their heart may faint, and their ruins be multiplied: ah! it is made bright, it is wrapped up for the

slaughter. cat ruinas in omnibus portis eorum dedi conturbationem gla-21:16 Go thee one way or other, either on the right hand, dii acuti et limati ad fulgendum amicti ad caedem 21:16 ex-

or on the left, whithersoever thy face is set.

acuere uade ad dextram siue ad sinistram quocumque faciei

21:17 I will also smite mine hands together, and I will tuae est appetitus 21:17 quin et ego plaudam manu ad ma-

cause my fury to rest: I the LORD have said it.

num et implebo indignationem meam ego Dominus locutus

21:18 The word of the LORD came unto me again, saying, sum 21:18 et factus est sermo Domini ad me dicens 21:19

21:19 Also, thou son of man, appoint thee two ways, that the sword of the king of Babylon may come: both twain shall come forth out of one land: and choose thou a place, choose it at the head of the way to the city.

et tu fili hominis pone tibi duas uias ut ueniat gladius regis Babylonis de terra una egredientur ambo et manu capiet con-

21:20 Appoint a way, that the sword may come to Rabbath iecturam in capite uiae ciuitatis coniciet 21:20 uiam pones

of the Ammonites, and to Judah in Jerusalem the defenced.

ut ueniat gladius ad Rabbath filiorum Ammon et ad Iudam

21:21 For the king of Babylon stood at the parting of in Hierusalem munitissimam 21:21 stetit enim rex Babylo-

the way, at the head of the two ways, to use divination: he made his arrows bright, he consulted with images, he

looked in the liver. nis in biuio in capite duarum uiarum diuinationem quaerens 21:22 At his right hand was the divination for Jerusalem, commiscens sagittas interrogauit idola exta consuluit 21:22

to appoint captains, to open the mouth in the slaughter, to lift up the voice with shouting, to appoint battering rams

against the gates, to cast a mount, and to build a fort. ad dextram eius facta est diuinatio super Hierusalem ut ponat arietes ut aperiat os in caede ut eleuet uocem in ululatu ut ponat arietes contra portas ut conportet aggerem ut aedificet

21:23 And it shall be unto them as a false divination in munitiones 21:23 eritque quasi consulens frustra oraculum

their sight, to them that have sworn oaths: but he will call to remembrance the iniquity, that they may be taken.

in oculis eorum et sabbatorum otium imitans ipse autem re-

21:24 Therefore thus saith the Lord GOD; Because ye cordabitur iniquitatis ad capiendum 21:24 idcirco haec dicit

have made your iniquity to be remembered, in that your transgressions are discovered, so that in all your doings your sins do appear; because, I say, that ye are come to remembrance, ye shall be taken with the hand.

Dominus Deus pro eo quod recordati estis iniquitatis uestrae et reuelastis praeuaricationes uestras et apparuerunt peccata uestra in omnibus cogitationibus uestris pro eo inquam quod

21:25 And thou, profane wicked prince of Israel, whose recordati estis manu capiemini 21:25 tu autem profane im-

day is come, when iniquity shall have an end,

pie dux Israhel cuius uenit dies in tempore iniquitatis prae-

21:26 Thus saith the Lord GOD; Remove the diadem, and finita 21:26 haec dicit Dominus Deus aufer cidarim tolle

take off the crown: this shall not be the same: exalt him that is low, and abase him that is high.

coronam nonne haec est quae humilem subleuauit et subli-

21:27 I will overturn, overturn, overturn, it: and it shall be mem humiliauit 21:27 iniquitatem iniquitatem iniquitatem no more, until he come whose right it is; and I will give it him.

ponam eam et hoc nunc factum est donec ueniret cuius est

21:28 And thou, son of man, prophesy and say, Thus saith iudicium et tradam ei 21:28 et tu fili hominis propheta et dic

the Lord GOD concerning the Ammonites, and concerning their reproach; even say thou, The sword, the sword is drawn: for the slaughter it is furbished, to consume because of the glittering:

haec dicit Dominus Deus ad filios Ammon et ad obprobrium eorum et dices mucro mucro euaginate ad occidendum li-

Prophetia Ezechielis

mate ut interficias et fulgeas 21:29 cum tibi uiderentur uana 21:29 Whiles they see vanity unto thee, whiles they divine

a lie unto thee, to bring thee upon the necks of them that are slain, of the wicked, whose day is come, when their

et diuinarentur mendacia ut dareris super colla uulneratorum iniquity shall have an end. impiorum quorum uenit dies in tempore iniquitatis praefinita

21:30 reuertere ad uaginam tuam in loco in quo creatus es in 21:30 Shall I cause it to return into his sheath? I will judge thee in the place where thou wast created, in the land of thy nativity.

terra natiuitatis tuae iudicabo te 21:31 et effundam super te

21:31 And I will pour out mine indignation upon thee, I will blow against thee in the fire of my wrath, and deliver thee into the hand of brutish men, and skilful to destroy.

indignationem meam in igne furoris mei sufflabo in te daboque te in manus hominum insipientium et fabricantium interitum 21:32 igni eris cibus sanguis tuus erit in medio terrae 21:32 Thou shalt be for fuel to the fire; thy blood shall be

in the midst of the land; thou shalt be no more remembered: for I the LORD have spoken it.

obliuioni traderis quia ego Dominus locutus sum

22:1 et factum est uerbum Domini ad me dicens 22:2 et 22:1 Moreover the word of the LORD came unto me, say-

ing,

22:2 Now, thou son of man, wilt thou judge, wilt thou judge the bloody city? yea, thou shalt shew her all her abominations.

tu fili hominis num iudicas num iudicas ciuitatem sanguinum 22:3 et ostendes ei omnes abominationes suas et di- 22:3 Then say thou, Thus saith the Lord GOD, The city

sheddeth blood in the midst of it, that her time may come, and maketh idols against herself to defile herself.

ces haec dicit Dominus Deus ciuitas effundens sanguinem in medio sui ut ueniat tempus eius et quae fecit idola contra semet ipsam ut pollueretur 22:4 in sanguine tuo qui a te 22:4 Thou art become guilty in thy blood that thou hast

shed; and hast defiled thyself in thine idols which thou hast made; and thou hast caused thy days to draw near, and art come even unto thy years: therefore have I made thee a reproach unto the heathen, and a mocking to all countries.

effusus est deliquisti et in idolis tuis quae fecisti polluta es et adpropinquare fecisti dies tuos et adduxisti tempus annorum tuorum propterea dedi te obprobrium gentibus et inrisionem uniuersis terris 22:5 quae iuxta sunt et quae procul 22:5 Those that be near, and those that be far from thee,

shall mock thee, which art infamous and much vexed.

a te triumphabunt de te sordida nobilis grandis interitu 22:6 22:6 Behold, the princes of Israel, every one were in thee

to their power to shed blood.

ecce principes Israhel singuli in brachio suo fuerunt in te ad effundendum sanguinem 22:7 patrem et matrem contume- 22:7 In thee have they set light by father and mother: in

the midst of thee have they dealt by oppression with the stranger: in thee have they vexed the fatherless and the

liis adfecerunt in te aduenam calumniati sunt in medio tui widow. pupillum et uiduam contristauerunt apud te 22:8 sanctuaria 22:8 Thou hast despised mine holy things, and hast pro-

faned my sabbaths.

mea spreuistis et sabbata mea polluistis 22:9 uiri detracto- 22:9 In thee are men that carry tales to shed blood: and in

thee they eat upon the mountains: in the midst of thee they commit lewdness.

res fuerunt in te ad effundendum sanguinem et super montes comederunt in te scelus operati sunt in medio tui 22:10 22:10 In thee have they discovered their fathers’ naked-

ness: in thee have they humbled her that was set apart for pollution.

uerecundiora patris discoperuerunt in te inmunditiam menstruatae humiliauerunt in te 22:11 et unusquisque in uxo- 22:11 And one hath committed abomination with his

neighbour’s wife; and another hath lewdly defiled his daughter in law; and another in thee hath humbled his sis-

rem proximi sui operatus est abominationem et socer nurum ter, his father’s daughter. suam polluit nefarie frater sororem suam filiam patris sui

oppressit in te 22:12 munera acceperunt apud te ad effun- 22:12 In thee have they taken gifts to shed blood; thou hast taken usury and increase, and thou hast greedily gained of thy neighbours by extortion, and hast forgotten me, saith

dendum sanguinem usuram et superabundantiam accepisti et the Lord GOD.

Biblia Sacra Vulgata

auare proximos tuos calumniabaris meique oblita es ait Do-

22:13 Behold, therefore I have smitten mine hand at thy minus Deus 22:13 ecce conplosi manus meas super auaridishonest gain which thou hast made, and at thy blood which hath been in the midst of thee.

tiam tuam quam fecisti et super sanguinem qui effusus est in

22:14 Can thine heart endure, or can thine hands be strong, medio tui 22:14 numquid sustinebit cor tuum aut praeualein the days that I shall deal with thee? I the LORD have spoken it, and will do it.

bunt manus tuae in diebus quos ego faciam tibi ego Dominus

22:15 And I will scatter thee among the heathen, and dis- locutus sum et faciam 22:15 et dispergam te in nationes et

perse thee in the countries, and will consume thy filthiness out of thee.

uentilabo te in terras et deficere faciam inmunditiam tuam a

22:16 And thou shalt take thine inheritance in thyself in te 22:16 et possidebo te in conspectu gentium et scies quia

the sight of the heathen, and thou shalt know that I am the LORD.

ego Dominus 22:17 et factum est uerbum Domini ad me

22:17 And the word of the LORD came unto me, saying,

22:18 Son of man, the house of Israel is to me become dicens 22:18 fili hominis uersa est mihi domus Israhel in

dross: all they are brass, and tin, and iron, and lead, in the midst of the furnace; they are even the dross of silver.

scoriam omnes isti aes et stagnum et ferrum et plumbum

22:19 Therefore thus saith the Lord GOD; Because ye are in medio fornacis scoria argenti facti sunt 22:19 propterea

all become dross, behold, therefore I will gather you into the midst of Jerusalem.

haec dicit Dominus Deus eo quod uersi estis omnes in scoriam propterea ecce ego congregabo uos in medium Hieru-

22:20 As they gather silver, and brass, and iron, and lead, salem 22:20 congregatione argenti et aeris et ferri et stagni

and tin, into the midst of the furnace, to blow the fire upon it, to melt it; so will I gather you in mine anger and in my

fury, and I will leave you there, and melt you. et plumbi in medium fornacis ut succendam in eam ignem

ad conflandum sic congregabo in furore meo et in ira mea

22:21 Yea, I will gather you, and blow upon you in the fire et requiescam et conflabo uos 22:21 et congregabo uos et of my wrath, and ye shall be melted in the midst therof.

succendam uos in igne furoris mei et conflabimini in medio

22:22 As silver is melted in the midst of the furnace, so eius 22:22 ut conflatur argentum in medio fornacis sic eritis

shall ye be melted in the midst thereof; and ye shall know that I the LORD have poured out my fury upon you.

in medio eius et scietis quia ego Dominus effuderim indigna-

22:23 And the word of the LORD came unto me, saying, tionem meam super uos 22:23 et factum est uerbum Domini 22:24 Son of man, say unto her, Thou art the land that is ad me dicens 22:24 fili hominis dic ei tu es terra inmunda

not cleansed, nor rained upon in the day of indignation.

22:25 There is a conspiracy of her prophets in the midst et non conpluta in die furoris 22:25 coniuratio prophetarum

thereof, like a roaring lion ravening the prey; they have devoured souls; they have taken the treasure and precious things; they have made her many widows in the midst thereof.

in medio eius sicut leo rugiens capiensque praedam animam deuorauerunt opes et pretium acceperunt uiduas eius multi-

22:26 Her priests have violated my law, and have profaned plicauerunt in medio illius 22:26 sacerdotes eius contemp-

mine holy things: they have put no difference between the holy and profane, neither have they shewed difference between the unclean and the clean, and have hid their eyes from my sabbaths, and I am profaned among them.

serunt legem meam et polluerunt sanctuaria mea inter sanctum et profanum non habuere distantiam et inter pollutum et mundum non intellexerunt et a sabbatis meis auerterunt ocu-

22:27 Her princes in the midst thereof are like wolves los suos et coinquinabar in medio eorum 22:27 principes

ravening the prey, to shed blood, and to destroy souls, to get dishonest gain.

eius in medio illius quasi lupi rapientes praedam ad effundendum sanguinem et perdendas animas et auare sectanda

Prophetia Ezechielis

lucra 22:28 prophetae autem eius liniebant eos absque tem- 22:28 And her prophets have daubed them with untem-

pered morter, seeing vanity, and divining lies unto them, saying, Thus saith the Lord GOD, when the LORD hath

peramento uidentes uana et diuinantes eis mendacium dicen- not spoken. tes haec dicit Dominus Deus cum Dominus non sit locutus 22:29 populi terrae calumniabantur calumniam et rapiebant 22:29 The people of the land have used oppression, and

exercised robbery, and have vexed the poor and needy: yea, they have oppressed the stranger wrongfully.

uiolenter egenum et pauperem adfligebant et aduenam opprimebant calumnia absque iudicio 22:30 et quaesiui de eis 22:30 And I sought for a man among them, that should

make up the hedge, and stand in the gap before me for the land, that I should not destroy it: but I found none.

uirum qui interponeret sepem et staret oppositus contra me pro terra ne dissiparem eam et non inueni 22:31 et effudi 22:31 Therefore have I poured out mine indignation upon

them; I have consumed them with the fire of my wrath: their own way have I recompensed upon their heads, saith

super eos indignationem meam in igne irae meae consumpsi the Lord GOD. eos uiam eorum in caput eorum reddidi ait Dominus Deus

23:1 et factus est sermo Domini ad me dicens 23:2 fili ho- 23:1 The word of the LORD came again unto me, saying,

23:2 Son of man, there were two women, the daughters of one mother:

minis duae mulieres filiae matris unius fuerunt 23:3 et for-

23:3 And they committed whoredoms in Egypt; they committed whoredoms in their youth: there were their breasts

nicatae sunt in Aegypto in adulescentia sua fornicatae sunt pressed, and there they bruised the teats of their virginity.

ibi subacta sunt ubera earum et fractae sunt mammae pubertatis earum 23:4 nomina autem earum Oolla maior et 23:4 And the names of them were Aholah the elder, and

Aholibah her sister: and they were mine, and they bare sons and daughters. Thus were their names; Samaria is

Ooliba soror eius et habui eas et pepererunt filios et filias Aholah, and Jerusalem Aholibah. porro earum nomina Samaria Oolla et Hierusalem Ooliba 23:5 fornicata est igitur Oolla super me et insaniuit in amato- 23:5 And Aholah played the harlot when she was mine;

and she doted on her lovers, on the Assyrians her neighbours,

res suos in Assyrios propinquantes 23:6 uestitos hyacintho

23:6 Which were clothed with blue, captains and rulers, all of them desirable young men, horsemen riding upon horses.

principes et magistratus iuuenes cupidinis uniuersos equites ascensores equorum 23:7 et dedit fornicationes suas su- 23:7 Thus she committed her whoredoms with them, with

all them that were the chosen men of Assyria, and with all on whom she doted: with all their idols she defiled herself.

per eos electos filios Assyriorum uniuersos et in omnibus in quos insaniuit in inmunditiis eorum polluta est 23:8 insu- 23:8 Neither left she her whoredoms brought from Egypt:

for in her youth they lay with her, and they bruised the breasts of her virginity, and poured their whoredom upon

per et fornicationes suas quas habuerat in Aegypto non reli- her. quit nam et illi dormierant cum ea in adulescentia eius et illi confregerant ubera pubertatis eius et effuderant fornicatio-

nem suam super eam 23:9 propterea tradidi eam in manu 23:9 Wherefore I have delivered her into the hand of her lovers, into the hand of the Assyrians, upon whom she doted.

amatorum suorum in manus filiorum Assur super quorum insaniuit libidinem 23:10 ipsi discoperuerunt ignominiam 23:10 These discovered her nakedness: they took her sons

and her daughters, and slew her with the sword: and she became famous among women; for they had executed

eius filios et filias illius tulerunt et ipsam occiderunt gladio judgment upon her. et factae sunt famosae mulieres et iudicia perpetrarunt in ea 23:11 quod cum uidisset soror eius Ooliba plus quam illa in- 23:11 And when her sister Aholibah saw this, she was

more corrupt in her inordinate love than she, and in her whoredoms more than her sister in her whoredoms.

saniuit libidine et fornicationem suam super fornicationem

Biblia Sacra Vulgata

23:12 She doted upon the Assyrians her neighbours, cap- sororis suae 23:12 ad filios Assyriorum praebuit inpudenter

tains and rulers clothed most gorgeously, horsemen riding upon horses, all of them desirable young men.

ducibus et magistratibus ad se uenientibus indutis ueste uaria equitibus qui uectabantur equis et adulescentibus forma

23:13 Then I saw that she was defiled, that they took both cunctis egregia 23:13 et uidi quod polluta esset uia una am-

one way,

23:14 And that she increased her whoredoms: for when barum 23:14 et auxit fornicationes suas cumque uidisset ui-

she saw men pourtrayed upon the wall, the images of the Chaldeans pourtrayed with vermilion,

ros depictos in pariete imagines Chaldeorum expressas colo-

23:15 Girded with girdles upon their loins, exceeding in ribus 23:15 et accinctos balteis renes et tiaras tinctas in ca-

dyed attire upon their heads, all of them princes to look to, after the manner of the Babylonians of Chaldea, the land

of their nativity: pitibus eorum formam ducum omnium similitudinem filio-23:16 And as soon as she saw them with her eyes, she rum Babylonis terraeque Chaldeorum in qua orti sunt 23:16

doted upon them, and sent messengers unto them into Chaldea.

et insaniuit super eos concupiscentia oculorum suorum et

23:17 And the Babylonians came to her into the bed of misit nuntios ad eos in Chaldeam 23:17 cumque uenissent

love, and they defiled her with their whoredom, and she was polluted with them, and her mind was alienated from

them. ad eam filii Babylonis ad cubile mammarum polluerunt eam

stupris suis et polluta est ab eis et saturata est anima eius ab

23:18 So she discovered her whoredoms, and discovered illis 23:18 denudauit quoque fornicationes suas et discope-

her nakedness: then my mind was alienated from her, like as my mind was alienated from her sister.

ruit ignominiam suam et recessit anima mea ab ea sicut re-

23:19 Yet she multiplied her whoredoms, in calling to re- cesserat anima mea a sorore eius 23:19 multiplicauit enim

membrance the days of her youth, wherein she had played the harlot in the land of Egypt.

fornicationes suas recordans dies adulescentiae suae quibus

23:20 For she doted upon their paramours, whose flesh is fornicata est in terra Aegypti 23:20 et insaniuit libidine su-

as the flesh of asses, and whose issue is like the issue of horses.

per concubitu eorum quorum carnes sunt ut carnes asinorum

23:21 Thus thou calledst to remembrance the lewdness of et sicut fluxus equorum fluxus eorum 23:21 et uisitasti sce-

thy youth, in bruising thy teats by the Egyptians for the paps of thy youth.

lus adulescentiae tuae quando subacta sunt in Aegypto ubera

23:22 Therefore, O Aholibah, thus saith the Lord GOD; tua et confractae mammae pubertatis tuae 23:22 propterea

Behold, I will raise up thy lovers against thee, from whom thy mind is alienated, and I will bring them against thee

on every side; Ooliba haec dicit Dominus Deus ecce ego suscitabo omnes

amatores tuos contra te de quibus satiata est anima tua et

23:23 The Babylonians, and all the Chaldeans, Pekod, and congregabo eos aduersum te in circuitu 23:23 filios Baby-Shoa, and Koa, and all the Assyrians with them: all of them desirable young men, captains and rulers, great lords

and renowned, all of them riding upon horses. lonis et uniuersos Chaldeos nobiles tyrannosque et principes omnes filios Assyriorum iuuenes forma egregia duces et magistratus uniuersos principes principum et nominatos as-

23:24 And they shall come against thee with chariots, censores equorum 23:24 et uenient super te instructi curru

wagons, and wheels, and with an assembly of people, which shall set against thee buckler and shield and helmet round about: and I will set judgment before them, and they shall judge thee according to their judgments.

et rota multitudo populorum lorica et clypeo et galea armabuntur contra te undique et dabo coram eis iudicium et iu-

23:25 And I will set my jealousy against thee, and they dicabunt te iudiciis suis 23:25 et ponam zelum meum in te

shall deal furiously with thee: they shall take away thy nose and thine ears; and thy remnant shall fall by the sword: they shall take thy sons and thy daughters; and thy residue shall be devoured by the fire.

quem exercent tecum in furore nasum tuum et aures tuas

Prophetia Ezechielis

praecident et quae remanserint gladio concident ipsi filios tuos et filias tuas capient et nouissimum tuum deuorabitur igni 23:26 et denudabunt te uestimentis tuis et tollent uasa 23:26 They shall also strip thee out of thy clothes, and take

away thy fair jewels.

gloriae tuae 23:27 et requiescere faciam scelus tuum de te et 23:27 Thus will I make thy lewdness to cease from thee,

and thy whoredom brought from the land of Egypt: so that thou shalt not lift up thine eyes unto them, nor remember

fornicationem tuam de terra Aegypti nec leuabis oculos tuos Egypt any more. ad eos et Aegypti non recordaberis amplius 23:28 quia haec 23:28 For thus saith the Lord GOD; Behold, I will deliver

thee into the hand of them whom thou hatest, into the hand of them from whom thy mind is alienated:

dicit Dominus Deus ecce ego tradam te in manu eorum quos odisti in manu de quibus satiata est anima tua 23:29 et agent 23:29 And they shall deal with thee hatefully, and shall

take away all thy labour, and shall leave thee naked and bare: and the nakedness of thy whoredoms shall be dis-

tecum in odio et tollent omnes labores tuos et dimittent te covered, both thy lewdness and thy whoredoms. nudam et ignominia plenam reuelabitur ignominia fornicationum tuarum scelus tuum et fornicationes tuae 23:30 fe- 23:30 I will do these things unto thee, because thou hast

gone a whoring after the heathen, and because thou art polluted with their idols.

cerunt haec tibi quia fornicata es post gentes inter quas polluta es in idolis eorum 23:31 in uia sororis tuae ambulasti 23:31 Thou hast walked in the way of thy sister; therefore

will I give her cup into thine hand.

et dabo calicem eius in manu tua 23:32 haec dicit Domi- 23:32 Thus saith the Lord GOD; Thou shalt drink of thy

sister’s cup deep and large: thou shalt be laughed to scorn and had in derision; it containeth much.

nus Deus calicem sororis tuae bibes profundum et latum eris in derisum et in subsannationem quae es capacissima 23:33 23:33 Thou shalt be filled with drunkenness and sorrow,

with the cup of astonishment and desolation, with the cup of thy sister Samaria.

ebrietate et dolore repleberis calice maeroris et tristitiae calice sororis tuae Samariae 23:34 et bibes illum et epotabis 23:34 Thou shalt even drink it and suck it out, and thou

shalt break the sherds thereof, and pluck off thine own breasts: for I have spoken it, saith the Lord GOD.

usque ad feces et fragmenta eius deuorabis et ubera tua lacerabis quia ego locutus sum ait Dominus Deus 23:35 propte- 23:35 Therefore thus saith the Lord GOD; Because thou

hast forgotten me, and cast me behind thy back, therefore bear thou also thy lewdness and thy whoredoms.

rea haec dicit Dominus Deus quia oblita es mei et proiecisti me post corpus tuum tu quoque porta scelus tuum et fornicationes tuas 23:36 et ait Dominus ad me dicens fili ho- 23:36 The LORD said moreover unto me; Son of man,

wilt thou judge Aholah and Aholibah? yea, declare unto them their abominations;

minis numquid iudicas Oollam et Oolibam et adnuntias eis scelera earum 23:37 quia adulterae sunt et sanguis in ma- 23:37 That they have committed adultery, and blood is in

their hands, and with their idols have they committed adultery, and have also caused their sons, whom they bare unto

nibus earum et cum idolis suis fornicatae sunt insuper et fi- me, to pass for them through the fire, to devour them. lios suos quos genuerunt mihi obtulerunt eis ad deuorandum 23:38 sed et hoc fecerunt mihi polluerunt sanctuarium meum 23:38 Moreover this they have done unto me: they have

defiled my sanctuary in the same day, and have profaned my sabbaths.

in die illa et sabbata mea profanauerunt 23:39 cumque im-

23:39 For when they had slain their children to their idols, then they came the same day into my sanctuary to profane it; and, lo, thus have they done in the midst of mine house.

molarent filios suos idolis suis et ingrederentur sanctuarium meum in die illa ut polluerent illud etiam haec fecerunt in

medio domus meae 23:40 miserunt ad uiros uenientes de 23:40 And furthermore, that ye have sent for men to come from far, unto whom a messenger was sent; and, lo, they came: for whom thou didst wash thyself, paintedst thy

longe ad quos nuntium miserant itaque ecce uenerunt qui- eyes, and deckedst thyself with ornaments,

Biblia Sacra Vulgata

bus te lauisti et circumleuisti stibio oculos tuos et ornata es

23:41 And satest upon a stately bed, and a table prepared mundo muliebri 23:41 sedisti in lecto pulcherrimo et mensa

before it, whereupon thou hast set mine incense and mine oil.

ordinata est ante te thymiama meum et unguentum meum

23:42 And a voice of a multitude being at ease was with posuisti super eam 23:42 et uox multitudinis exultantis erat

her: and with the men of the common sort were brought Sabeans from the wilderness, which put bracelets upon

their hands, and beautiful crowns upon their heads. in ea et in uiris qui de multitudine hominum adducebantur et

ueniebant de deserto posuerunt armillas in manibus eorum et

23:43 Then said I unto her that was old in adulteries, Will coronas speciosas in capitibus eorum 23:43 et dixi ei quae

they now commit whoredoms with her, and she with them?

adtrita est in adulteriis nunc fornicabitur in fornicatione sua

23:44 Yet they went in unto her, as they go in unto a etiam haec 23:44 et ingressi sunt ad eam quasi ad mulie-

woman that playeth the harlot: so went they in unto Aholah and unto Aholibah, the lewd women.

rem meretricem sic ingrediebantur ad Oollam et ad Oolibam

23:45 And the righteous men, they shall judge them after mulieres nefarias 23:45 uiri ergo iusti sunt hii iudicabunt

the manner of adulteresses, and after the manner of women that shed blood; because they are adulteresses, and blood

is in their hands. eas iudicio adulterarum et iudicio effundentium sanguinem 23:46 For thus saith the Lord GOD; I will bring up a com- quia adulterae sunt et sanguis in manibus earum 23:46 haec

pany upon them, and will give them to be removed and spoiled.

enim dicit Dominus Deus adduc ad eas multitudinem et trade

23:47 And the company shall stone them with stones, and eas in tumultum et in rapinam 23:47 et lapidentur lapidibus

dispatch them with their swords; they shall slay their sons and their daughters, and burn up their houses with fire.

populorum et confodiantur gladiis eorum filios et filias ea-

23:48 Thus will I cause lewdness to cease out of the land, rum interficient et domos earum igne succendent 23:48 et that all women may be taught not to do after your lewdness.

auferam scelus de terra et discent omnes mulieres ne faciant

23:49 And they shall recompense your lewdness upon you, secundum scelus earum 23:49 et dabunt scelus uestrum su-

and ye shall bear the sins of your idols: and ye shall know that I am the Lord GOD.

per uos et peccata idolorum uestrorum portabitis et scietis quia ego Dominus Deus

24:1 Again in the ninth year, in the tenth month, in the 24:1 et factum est uerbum Domini ad me in anno nono in tenth day of the month, the word of the LORD came unto me, saying,

mense decimo decima mensis dicens 24:2 fili hominis sc-

24:2 Son of man, write thee the name of the day, even of this same day: the king of Babylon set himself against Jerusalem this same day.

ribe tibi nomen diei huius in qua confirmatus est rex Baby-

24:3 And utter a parable unto the rebellious house, and say lonis aduersum Hierusalem hodie 24:3 et dices per prouerunto them, Thus saith the Lord GOD; Set on a pot, set it on, and also pour water into it:

bium ad domum inritatricem parabolam et loqueris ad eos haec dicit Dominus Deus pone ollam pone inquam et mitte in

24:4 Gather the pieces thereof into it, even every good ea aquam 24:4 congere frusta eius in ea omnem partem bopiece, the thigh, and the shoulder; fill it with the choice bones.

nam femur et armum electa et ossibus plena 24:5 pinguissi-

24:5 Take the choice of the flock, and burn also the bones under it, and make it boil well, and let them seethe the bones of it therein.

mum pecus adsume conpone quoque struices ossuum sub ea efferbuit coctio eius et discocta sunt ossa illius in medio eius

24:6 Wherefore thus saith the Lord GOD; Woe to the 24:6 propterea haec dicit Dominus Deus uae ciuitati sangui-

bloody city, to the pot whose scum is therein, and whose scum is not gone out of it! bring it out piece by piece; let

no lot fall upon it. num ollae cuius rubigo in ea est et rubigo eius non exiuit

Prophetia Ezechielis

de ea per partes et per partes suas eice eam non cecidit super eam sors 24:7 sanguis enim eius in medio eius est super 24:7 For her blood is in the midst of her; she set it upon

the top of a rock; she poured it not upon the ground, to cover it with dust;

limpidissimam petram effudit illum non effudit illum super terram ut possit operiri puluere 24:8 ut superducerem indi- 24:8 That it might cause fury to come up to take

vengeance; I have set her blood upon the top of a rock, that it should not be covered.

gnationem meam et uindicta ulciscerer dedi sanguinem eius super petram limpidissimam ne operiretur 24:9 propterea 24:9 Therefore thus saith the Lord GOD; Woe to the

bloody city! I will even make the pile for fire great.

haec dicit Dominus Deus uae ciuitati sanguinum cuius ego grandem faciam pyram 24:10 congere ossa quae igne suc- 24:10 Heap on wood, kindle the fire, consume the flesh,

and spice it well, and let the bones be burned.

cendam consumentur carnes et concoquetur uniuersa conpositio et ossa tabescent 24:11 pone quoque eam super pru- 24:11 Then set it empty upon the coals thereof, that the

brass of it may be hot, and may burn, and that the filthiness of it may be molten in it, that the scum of it may be

nas uacuam ut incalescat et liquefiat aes eius et confletur in consumed. medio eius inquinamentum eius et consumatur rubigo eius 24:12 multo labore sudatum est et non exibit de ea nimia ru- 24:12 She hath wearied herself with lies, and her great

scum went not forth out of her: her scum shall be in the fire.

bigo eius neque per ignem 24:13 inmunditia tua execrabi-

24:13 In thy filthiness is lewdness: because I have purged thee, and thou wast not purged, thou shalt not be purged from thy filthiness any more, till I have caused my fury to rest upon thee.

lis quia mundare te uolui et non es mundata a sordibus tuis sed nec mundaberis prius donec quiescere faciam indignationem meam in te 24:14 ego Dominus locutus sum uenit 24:14 I the LORD have spoken it: it shall come to pass,

and I will do it; I will not go back, neither will I spare, neither will I repent; according to thy ways, and according

et faciam non transeam nec parcam nec placabor iuxta uias to thy doings, shall they judge thee, saith the Lord GOD. tuas et iuxta adinuentiones tuas iudicaui te dicit Dominus 24:15 et factum est uerbum Domini ad me dicens 24:16 fili 24:15 Also the word of the LORD came unto me, saying,

24:16 Son of man, behold, I take away from thee the desire of thine eyes with a stroke: yet neither shalt thou mourn nor weep, neither shall thy tears run down.

hominis ecce ego tollo a te desiderabile oculorum tuorum in plaga et non planges neque plorabis neque fluent lacrimae tuae 24:17 ingemesce tacens mortuorum luctum non facies 24:17 Forbear to cry, make no mourning for the dead, bind

the tire of thine head upon thee, and put on thy shoes upon thy feet, and cover not thy lips, and eat not the bread of

corona tua circumligata sit tibi et calciamenta tua erunt in men. pedibus tuis nec amictu ora uelabis nec cibos lugentium comedes 24:18 locutus sum ergo ad populum mane et mortua 24:18 So I spake unto the people in the morning: and at

even my wife died; and I did in the morning as I was commanded.

est uxor mea uesperi fecique mane sicut praeceperat mihi 24:19 et dixit ad me populus quare non indicas nobis quid ista 24:19 And the people said unto me, Wilt thou not tell us

what these things are to us, that thou doest so?

significent quae tu facis 24:20 et dixi ad eos sermo Domini 24:20 Then I answered them, The word of the LORD came

unto me, saying,

factus est ad me dicens 24:21 loquere domui Israhel haec 24:21 Speak unto the house of Israel, Thus saith the Lord

GOD; Behold, I will profane my sanctuary, the excellency of your strength, the desire of your eyes, and that which your soul pitieth; and your sons and your daughters whom ye have left shall fall by the sword.

dicit Dominus Deus ecce ego polluam sanctuarium meum superbiam imperii uestri et desiderabile oculorum uestrorum et super quo pauet anima uestra et filii uestri et filiae quas re-

Biblia Sacra Vulgata

24:22 And ye shall do as I have done: ye shall not cover liquistis gladio cadent 24:22 et facietis sicut feci ora amictu

your lips, nor eat the bread of men.

24:23 And your tires shall be upon your heads, and your non uelabitis et cibos lugentium non comedetis 24:23 coro-

shoes upon your feet: ye shall not mourn nor weep; but ye shall pine away for your iniquities, and mourn one toward

another. nas habebitis in capitibus uestris et calciamenta in pedibus

non plangetis neque flebitis sed tabescetis in iniquitatibus

24:24 Thus Ezekiel is unto you a sign: according to all uestris et unusquisque gemet ad fratrem suum 24:24 erit-

that he hath done shall ye do: and when this cometh, ye shall know that I am the Lord GOD.

que Hiezecihel uobis in portentum iuxta omnia quae fecit facietis cum uenerit istud et scietis quia ego Dominus Deus

24:25 Also, thou son of man, shall it not be in the day 24:25 et tu fili hominis ecce in die quo tollam ab eis fortitu-

when I take from them their strength, the joy of their glory, the desire of their eyes, and that whereupon they set their

minds, their sons and their daughters, dinem eorum et gaudium dignitatis et desiderium oculorum

eorum super quo requiescunt animae eorum filios et filias eo-

24:26 That he that escapeth in that day shall come unto rum 24:26 in die illa cum uenerit fugiens ad te ut adnuntiet

thee, to cause thee to hear it with thine ears?

24:27 In that day shall thy mouth be opened to him which tibi 24:27 in die inquam illa aperietur os tuum cum eo qui

is escaped, and thou shalt speak, and be no more dumb: and thou shalt be a sign unto them; and they shall know

that I am the LORD. fugit et loqueris et non silebis ultra erisque eis in portentum et scietis quia ego Dominus

25:1 The word of the LORD came again unto me, saying, 25:1 et factus est sermo Domini ad me dicens 25:2 fili ho-

25:2 Son of man, set thy face against the Ammonites, and prophesy against them;

minis pone faciem tuam contra filios Ammon et prophetabis

25:3 And say unto the Ammonites, Hear the word of de eis 25:3 et dices filiis Ammon audite uerbum Domini

the Lord GOD; Thus saith the Lord GOD; Because thou saidst, Aha, against my sanctuary, when it was profaned; and against the land of Israel, when it was desolate; and against the house of Judah, when they went into captivity;

Dei haec dicit Dominus Deus pro eo quod dixisti euge euge super sanctuarium meum quia pollutum est et super terram Israhel quoniam desolata est et super domum Iuda quoniam

25:4 Behold, therefore I will deliver thee to the men of the ducti sunt in captiuitatem 25:4 idcirco ego tradam te filiis

east for a possession, and they shall set their palaces in thee, and make their dwellings in thee: they shall eat thy

fruit, and they shall drink thy milk. orientalibus in hereditatem et conlocabunt caulas suas in te

et ponent in te tentoria sua ipsi comedent fruges tuas et ipsi

25:5 And I will make Rabbah a stable for camels, and the bibent lac tuum 25:5 daboque Rabbath in habitaculum ca-

Ammonites a couching place for flocks: and ye shall know that I am the LORD.

melorum et filios Ammon in cubile pecorum et scietis quia

25:6 For thus saith the Lord GOD; Because thou hast ego Dominus 25:6 quia haec dicit Dominus Deus pro eo

clapped thine hands, and stamped with the feet, and rejoiced in heart with all thy despite against the land of Is-

rael; quod plausisti manu et percussisti pede et gauisa es ex toto 25:7 Behold, therefore I will stretch out mine hand upon affectu super terram Israhel 25:7 idcirco ecce ego extendam

thee, and will deliver thee for a spoil to the heathen; and I will cut thee off from the people, and I will cause thee to perish out of the countries: I will destroy thee; and thou shalt know that I am the LORD.

manum meam super te et tradam te in direptionem gentium et interficiam te de populis et perdam de terris et conteram

25:8 Thus saith the Lord GOD; Because that Moab and et scies quia ego Dominus 25:8 haec dicit Dominus Deus

Seir do say, Behold, the house of Judah is like unto all the heathen;

pro eo quod dixerunt Moab et Seir ecce sicut omnes gentes

25:9 Therefore, behold, I will open the side of Moab domus Iuda 25:9 idcirco ecce ego aperiam umerum Moab

from the cities, from his cities which are on his frontiers, the glory of the country, Bethjeshimoth, Baalmeon, and Kiriathaim,

Prophetia Ezechielis

de ciuitatibus de ciuitatibus inquam eius et de finibus eius inclitas terrae Bethiesimoth et Beelmeon et Cariathaim 25:10 25:10 Unto the men of the east with the Ammonites, and

will give them in possession, that the Ammonites may not be remembered among the nations.

filiis orientis cum filiis Ammon et dabo eam in hereditatem ut non sit memoria ultra filiorum Ammon in gentibus 25:11 25:11 And I will execute judgments upon Moab; and they

shall know that I am the LORD.

et in Moab faciam iudicia et scient quia ego Dominus 25:12 25:12 Thus saith the Lord GOD; Because that Edom hath

dealt against the house of Judah by taking vengeance, and hath greatly offended, and revenged himself upon them;

haec dicit Dominus Deus pro eo quod fecit Idumea ultionem ut se uindicaret de filiis Iuda peccauitque delinquens et uindictam expetiuit de eis 25:13 idcirco haec dicit Dominus 25:13 Therefore thus saith the Lord GOD; I will also

stretch out mine hand upon Edom, and will cut off man and beast from it; and I will make it desolate from Teman;

Deus extendam manum meam super Idumeam et auferam de and they of Dedan shall fall by the sword. ea hominem et iumentum et faciam eam desertum ab austro et qui sunt in Daedan gladio cadent 25:14 et dabo ultionem 25:14 And I will lay my vengeance upon Edom by the

hand of my people Israel: and they shall do in Edom according to mine anger and according to my fury; and they

meam super Idumeam per manum populi mei Israhel et fa- shall know my vengeance, saith the Lord GOD. cient in Edom iuxta iram meam et furorem meum et scient uindictam meam dicit Dominus Deus 25:15 haec dicit Do- 25:15 Thus saith the Lord GOD; Because the Philistines

have dealt by revenge, and have taken vengeance with a despiteful heart, to destroy it for the old hatred;

minus Deus pro eo quod fecerunt Palestini in uindictam et ulti se sunt toto animo interficientes et implentes inimicitias ueteres 25:16 propterea haec dicit Dominus Deus ecce 25:16 Therefore thus saith the Lord GOD; Behold, I will

stretch out mine hand upon the Philistines, and I will cut off the Cherethims, and destroy the remnant of the sea

ego extendam manum meam super Palestinos et interficiam coast. interfectores et perdam reliquias maritimae regionis 25:17 25:17 And I will execute great vengeance upon them with

furious rebukes; and they shall know that I am the LORD, when I shall lay my vengeance upon them.

faciamque in eis ultiones magnas arguens in furore et scient quia ego Dominus cum dedero uindictam meam super eos

26:1 et factum est in undecimo anno prima mensis factus 26:1 And it came to pass in the eleventh year, in the first

day of the month, that the word of the LORD came unto me, saying,

est sermo Domini ad me dicens 26:2 fili hominis pro eo

26:2 Son of man, because that Tyrus hath said against Jerusalem, Aha, she is broken that was the gates of the people: she is turned unto me: I shall be replenished, now she is laid waste:

quod dixit Tyrus de Hierusalem euge confractae sunt portae populorum conuersa est ad me implebor deserta est 26:3 26:3 Therefore thus saith the Lord GOD; Behold, I am

against thee, O Tyrus, and will cause many nations to come up against thee, as the sea causeth his waves to come

propterea haec dicit Dominus Deus ecce ego super te Tyre up. et ascendere faciam ad te gentes multas sicut ascendit mare fluctuans 26:4 et dissipabunt muros Tyri et destruent tur- 26:4 And they shall destroy the walls of Tyrus, and break

down her towers: I will also scrape her dust from her, and make her like the top of a rock.

res eius et radam puluerem eius de ea et dabo eam in limpidissimam petram 26:5 siccatio sagenarum erit in medio 26:5 It shall be a place for the spreading of nets in the

midst of the sea: for I have spoken it, saith the Lord GOD: and it shall become a spoil to the nations.

maris quia ego locutus sum ait Dominus Deus et erit in direptionem gentibus 26:6 filiae quoque eius quae sunt in agro 26:6 And her daughters which are in the field shall be slain

by the sword; and they shall know that I am the LORD.

gladio interficientur et scient quia ego Dominus 26:7 quia 26:7 For thus saith the Lord GOD; Behold, I will bring

upon Tyrus Nebuchadrezzar king of Babylon, a king of kings, from the north, with horses, and with chariots, and with horsemen, and companies, and much people.

Biblia Sacra Vulgata

haec dicit Dominus Deus ecce ego adducam ad Tyrum Nabuchodonosor regem Babylonis ab aquilone regem regum cum equis et curribus et equitibus et coetu populoque magno

26:8 He shall slay with the sword thy daughters in the 26:8 filias tuas quae sunt in agro gladio interficiet et circum-

field: and he shall make a fort against thee, and cast a mount against thee, and lift up the buckler against thee.

dabit te munitionibus et conportabit aggerem in gyro et le-

26:9 And he shall set engines of war against thy walls, and uabit contra te clypeum 26:9 et uineas et arietes temperabit

with his axes he shall break down thy towers.

26:10 By reason of the abundance of his horses their dust in muros tuos et turres tuas destruet in armatura sua 26:10

shall cover thee: thy walls shall shake at the noise of the horsemen, and of the wheels, and of the chariots, when he shall enter into thy gates, as men enter into a city wherein is made a breach.

inundatione equorum eius operiet te puluis eorum a sonitu equitum et rotarum et curruum mouebuntur muri tui dum ingressus fuerit portas tuas quasi per introitus urbis dissipa-

26:11 With the hoofs of his horses shall he tread down all tae 26:11 ungulis equorum suorum conculcabit omnes pla-

thy streets: he shall slay thy people by the sword, and thy strong garrisons shall go down to the ground.

teas tuas populum tuum gladio caedet et statuae tuae nobiles

26:12 And they shall make a spoil of thy riches, and make in terram corruent 26:12 uastabunt opes tuas diripient nego-

a prey of thy merchandise: and they shall break down thy walls, and destroy thy pleasant houses: and they shall lay thy stones and thy timber and thy dust in the midst of the water.

tiationes tuas et destruent muros tuos et domos tuas praeclaras subuertent et lapides tuos et ligna tua et puluerem tuum in

26:13 And I will cause the noise of thy songs to cease; and medio aquarum ponent 26:13 et quiescere faciam multitu-

the sound of thy harps shall be no more heard.

dinem canticorum tuorum et sonitus cithararum tuarum non

26:14 And I will make thee like the top of a rock: thou audietur amplius 26:14 et dabo te in limpidissimam petram

shalt be a place to spread nets upon; thou shalt be built no more: for I the LORD have spoken it, saith the Lord GOD.

siccatio sagenarum eris nec aedificaberis ultra quia ego locu-

26:15 Thus saith the Lord GOD to Tyrus; Shall not the tus sum dicit Dominus Deus 26:15 haec dicit Dominus De-

isles shake at the sound of thy fall, when the wounded cry, when the slaughter is made in the midst of thee?

us Tyro numquid non a sonitu ruinae tuae et gemitu interfectorum tuorum cum occisi fuerint in medio tui commouebun-

26:16 Then all the princes of the sea shall come down tur insulae 26:16 et descendent de sedibus suis omnes prin-

from their thrones, and lay away their robes, and put off their broidered garments: they shall clothe themselves with trembling; they shall sit upon the ground, and shall tremble at every moment, and be astonished at thee.

cipes maris et auferent exuuias suas et uestimenta sua uaria abicient et induentur stupore in terra sedebunt et adtoniti su-

26:17 And they shall take up a lamentation for thee, and per repentino casu tuo admirabuntur 26:17 et adsumentes

say to thee, How art thou destroyed, that wast inhabited of seafaring men, the renowned city, which wast strong in the sea, she and her inhabitants, which cause their terror to be on all that haunt it!

super te lamentum dicent tibi quomodo peristi quae habitas in mari urbs inclita quae fuisti fortis in mari cum habitato-

26:18 Now shall the isles tremble in the day of thy fall; ribus tuis quos formidabant uniuersi 26:18 nunc stupebunt

yea, the isles that are in the sea shall be troubled at thy departure.

naues in die pauoris tui et turbabuntur insulae in mari eo

26:19 For thus saith the Lord GOD; When I shall make quod nullus egrediatur ex te 26:19 quia haec dicit Dominus

thee a desolate city, like the cities that are not inhabited; when I shall bring up the deep upon thee, and great waters

shall cover thee; Deus cum dedero te urbem desolatam sicut ciuitates quae non habitantur et adduxero super te abyssum et operuerint te

Prophetia Ezechielis

aquae multae 26:20 et detraxero te cum his qui descendunt 26:20 When I shall bring thee down with them that de-

scend into the pit, with the people of old time, and shall set thee in the low parts of the earth, in places desolate of old, with them that go down to the pit, that thou be not inhabited; and I shall set glory in the land of the living;

in lacum ad populum sempiternum et conlocauero te in terra nouissima sicut solitudines ueteres cum his qui deducuntur in lacum ut non habiteris porro dedero gloriam in terra uiuentium 26:21 in nihilum rediiam te et non eris et requisita 26:21 I will make thee a terror, and thou shalt be no more:

though thou be sought for, yet shalt thou never be found again, saith the Lord GOD.

non inuenieris ultra in sempiternum dicit Dominus Deus

27:1 et factum est uerbum Domini ad me dicens 27:2 tu 27:1 The word of the LORD came again unto me, saying,

27:2 Now, thou son of man, take up a lamentation for Tyrus;

ergo fili hominis adsume super Tyrum lamentum 27:3 et di-

27:3 And say unto Tyrus, O thou that art situate at the entry of the sea, which art a merchant of the people for many isles, Thus saith the Lord GOD; O Tyrus, thou hast said, I am of perfect beauty.

ces Tyro quae habitat in introitu maris negotiationi populorum ad insulas multas haec dicit Dominus Deus o Tyre tu dixisti perfecti decoris ego sum 27:4 et in corde maris sita 27:4 Thy borders are in the midst of the seas, thy builders

have perfected thy beauty.

finitimi tui qui te aedificauerunt impleuerunt decorem tuum 27:5 abietibus de Sanir extruxerunt te cum omnibus tabula- 27:5 They have made all thy ship boards of fir trees of

Senir: they have taken cedars from Lebanon to make masts for thee.

tis maris cedrum de Libano tulerunt ut facerent tibi malum 27:6 quercus de Basan dolauerunt in remos tuos transtra tua 27:6 Of the oaks of Bashan have they made thine oars; the

company of the Ashurites have made thy benches of ivory, brought out of the isles of Chittim.

fecerunt tibi ex ebore indico et praetoriola de insulis Italiae 27:7 byssus uaria de Aegypto texta est tibi in uelum ut pone- 27:7 Fine linen with broidered work from Egypt was that

which thou spreadest forth to be thy sail; blue and purple from the isles of Elishah was that which covered thee.

retur in malo hyacinthus et purpura de insulis Elisa facta sunt operimentum tuum 27:8 habitatores Sidonis et Aradii fue- 27:8 The inhabitants of Zidon and Arvad were thy

mariners: thy wise men, O Tyrus, that were in thee, were thy pilots.

runt remiges tui sapientes tui Tyre facti sunt gubernatores tui 27:9 senes Bibli et prudentes eius habuerunt nautas ad minis- 27:9 The ancients of Gebal and the wise men thereof were

in thee thy calkers: all the ships of the sea with their mariners were in thee to occupy thy merchandise.

terium uariae supellectilis tuae omnes naues maris et nautae earum fuerunt in populo negotiationis tuae 27:10 Persae et 27:10 They of Persia and of Lud and of Phut were in thine

army, thy men of war: they hanged the shield and helmet in thee; they set forth thy comeliness.

Lydi et Lybies erant in exercitu tuo uiri bellatores tui clypeum et galeam suspenderunt in te pro ornatu tuo 27:11 fi- 27:11 The men of Arvad with thine army were upon thy

walls round about, and the Gammadims were in thy towers: they hanged their shields upon thy walls round about;

lii Aradii cum exercitu tuo erant super muros tuos in circuitu they have made thy beauty perfect. sed et Pigmei qui erant in turribus tuis faretras suas suspenderunt in muris tuis per gyrum ipsi conpleuerunt pulchritudinem tuam 27:12 Carthaginienses negotiatores tui a mul- 27:12 Tarshish was thy merchant by reason of the multi-

tude of all kind of riches; with silver, iron, tin, and lead, they traded in thy fairs.

titudine cunctarum diuitiarum argento ferro stagno plumboque repleuerunt nundinas tuas 27:13 Graecia Thubal et Mo- 27:13 Javan, Tubal, and Meshech, they were thy mer-

chants: they traded the persons of men and vessels of brass in thy market.

soch ipsi institores tui mancipia et uasa aerea adduxerunt

populo tuo 27:14 de domo Thogorma equos et equites et 27:14 They of the house of Togarmah traded in thy fairs with horses and horsemen and mules.

Biblia Sacra Vulgata

27:15 The men of Dedan were thy merchants; many isles mulos adduxerunt ad forum tuum 27:15 filii Dadan negotia-

were the merchandise of thine hand: they brought thee for a present horns of ivory and ebony.

tores tui insulae multae negotiatio manus tuae dentes ebur-

27:16 Syria was thy merchant by reason of the multitude neos et hebeninos commutauerunt in pretio tuo 27:16 Syrus

of the wares of thy making: they occupied in thy fairs with emeralds, purple, and broidered work, and fine linen, and

coral, and agate. negotiator tuus propter multitudinem operum tuorum gem-

mam purpuram et scutulata et byssum et sericum et chod-

27:17 Judah, and the land of Israel, they were thy mer- chod proposuerunt in mercatu tuo 27:17 Iuda et terra Isra-

chants: they traded in thy market wheat of Minnith, and Pannag, and honey, and oil, and balm.

hel ipsi institores tui in frumento primo balsamum et mel et

27:18 Damascus was thy merchant in the multitude of the oleum et resinam proposuerunt in nundinis tuis 27:18 Da-

wares of thy making, for the multitude of all riches; in the wine of Helbon, and white wool.

mascenus negotiator tuus in multitudine operum tuorum in multitudine diuersarum opum in uino pingui in lanis colo-

27:19 Dan also and Javan going to and fro occupied in thy ris optimi 27:19 Dan et Graecia et Mozel in nundinis tuis

fairs: bright iron, cassia, and calamus, were in thy market.

proposuerunt ferrum fabrefactum stacte et calamus in nego-

27:20 Dedan was thy merchant in precious clothes for tiatione tua 27:20 Dadan institores tui in tapetibus ad seden-

chariots.

27:21 Arabia, and all the princes of Kedar, they occupied dum 27:21 Arabia et uniuersi principes Cedar ipsi negotia-

with thee in lambs, and rams, and goats: in these were they thy merchants.

tores manus tuae cum agnis et arietibus et hedis uenerunt ad

27:22 The merchants of Sheba and Raamah, they were te negotiatores tui 27:22 uenditores Saba et Reema ipsi ne-

thy merchants: they occupied in thy fairs with chief of all spices, and with all precious stones, and gold.

gotiatores tui cum uniuersis primis aromatibus et lapide pre-

27:23 Haran, and Canneh, and Eden, the merchants of tioso et auro quod proposuerunt in mercatu tuo 27:23 Aran

Sheba, Asshur, and Chilmad, were thy merchants.

et Chenne et Eden negotiatores Saba Assur Chelmad uendi-

27:24 These were thy merchants in all sorts of things, in tores tui 27:24 ipsi negotiatores tui multifariam inuolucris

blue clothes, and broidered work, and in chests of rich apparel, bound with cords, and made of cedar, among thy

merchandise. hyacinthi et polymitorum gazarumque pretiosarum quae ob-

uolutae et adstrictae erant funibus cedros quoque habebant in

27:25 The ships of Tarshish did sing of thee in thy market: negotiationibus tuis 27:25 naues maris principes tuae in ne-

and thou wast replenished, and made very glorious in the midst of the seas.

gotiatione tua et repleta es et glorificata nimis in corde maris

27:26 Thy rowers have brought thee into great waters: the 27:26 in aquis multis adduxerunt te remiges tui uentus auster

east wind hath broken thee in the midst of the seas.

27:27 Thy riches, and thy fairs, thy merchandise, thy contriuit te in corde maris 27:27 diuitiae tuae et thesauri tui

mariners, and thy pilots, thy calkers, and the occupiers of thy merchandise, and all thy men of war, that are in thee, and in all thy company which is in the midst of thee, shall fall into the midst of the seas in the day of thy ruin.

et multiplex instrumentum tuum nautae tui et gubernatores tui qui tenebant supellectilem tuam et populo tuo praeerant uiri quoque bellatores tui qui erant in te cum uniuersa multitudine tua quae est in medio tui cadent in corde maris in

27:28 The suburbs shall shake at the sound of the cry of die ruinae tuae 27:28 a sonitu clamoris gubernatorum tuo-

thy pilots.

27:29 And all that handle the oar, the mariners, and all the rum conturbabuntur classes 27:29 et descendent de nauibus

pilots of the sea, shall come down from their ships, they shall stand upon the land;

suis omnes qui tenebant remum nautae et uniuersi guber-

Prophetia Ezechielis

natores maris in terra stabunt 27:30 et heiulabunt super te 27:30 And shall cause their voice to be heard against thee,

and shall cry bitterly, and shall cast up dust upon their heads, they shall wallow themselves in the ashes:

uoce magna et clamabunt amare et superiacient puluerem capitibus suis et cinere conspergentur 27:31 et radent super 27:31 And they shall make themselves utterly bald for

thee, and gird them with sackcloth, and they shall weep for thee with bitterness of heart and bitter wailing.

te caluitium et accingentur ciliciis et plorabunt te in amaritudine animae ploratu amarissimo 27:32 et adsument super te 27:32 And in their wailing they shall take up a lamentation

for thee, and lament over thee, saying, What city is like Tyrus, like the destroyed in the midst of the sea?

carmen lugubre et plangent te quae est ut Tyrus quae obmutuit in medio maris 27:33 quae in exitu negotiationum tua- 27:33 When thy wares went forth out of the seas, thou

filledst many people; thou didst enrich the kings of the earth with the multitude of thy riches and of thy merchan-

rum de mari implesti populos multos in multitudine diuitia- dise. rum tuarum et populorum tuorum ditasti reges terrae 27:34 27:34 In the time when thou shalt be broken by the seas

in the depths of the waters thy merchandise and all thy company in the midst of thee shall fall.

nunc contrita es a mari in profundis aquarum opes tuae et omnis multitudo tua quae erat in medio tui ceciderunt 27:35 27:35 All the inhabitants of the isles shall be astonished

at thee, and their kings shall be sore afraid, they shall be troubled in their countenance.

uniuersi habitatores insularum obstipuerunt super te et reges earum omnes tempestate perculsi mutauerunt uultus 27:36 27:36 The merchants among the people shall hiss at thee;

thou shalt be a terror, and never shalt be any more.

negotiatores populorum sibilauerunt super te ad nihilum deducta es et non eris usque in perpetuum

28:1 et factus est sermo Domini ad me dicens 28:2 fili 28:1 The word of the LORD came again unto me, saying,

28:2 Son of man, say unto the prince of Tyrus, Thus saith the Lord GOD; Because thine heart is lifted up, and thou hast said, I am a God, I sit in the seat of God, in the midst of the seas; yet thou art a man, and not God, though thou set thine heart as the heart of God:

hominis dic principi Tyri haec dicit Dominus Deus eo quod eleuatum est cor tuum et dixisti Deus ego sum et in cathedra Dei sedi in corde maris cum sis homo et non Deus et dedisti

cor tuum quasi cor Dei 28:3 ecce sapientior es tu Danihele 28:3 Behold, thou art wiser than Daniel; there is no secret

that they can hide from thee:

omne secretum non est absconditum a te 28:4 in sapientia 28:4 With thy wisdom and with thine understanding thou

hast gotten thee riches, and hast gotten gold and silver into thy treasures:

et prudentia tua fecisti tibi fortitudinem et adquisisti aurum et argentum in thesauris tuis 28:5 in multitudine sapientiae 28:5 By thy great wisdom and by thy traffick hast thou

increased thy riches, and thine heart is lifted up because of thy riches:

tuae et in negotiatione tua multiplicasti tibi fortitudinem et eleuatum est cor tuum in robore tuo 28:6 propterea haec di- 28:6 Therefore thus saith the Lord GOD; Because thou

hast set thine heart as the heart of God;

cit Dominus Deus eo quod eleuatum est cor tuum quasi cor

Dei 28:7 idcirco ecce ego adducam super te alienos robus- 28:7 Behold, therefore I will bring strangers upon thee, the terrible of the nations: and they shall draw their swords against the beauty of thy wisdom, and they shall defile thy

tissimos gentium et nudabunt gladios suos super pulchritu- brightness. dinem sapientiae tuae et polluent decorem tuum 28:8 inter- 28:8 They shall bring thee down to the pit, and thou shalt

die the deaths of them that are slain in the midst of the seas.

ficient et detrahent te et morieris interitu occisorum in corde

maris 28:9 numquid dicens loqueris Deus ego sum coram 28:9 Wilt thou yet say before him that slayeth thee, I am God? but thou shalt be a man, and no God, in the hand of him that slayeth thee.

interficientibus te cum sis homo et non Deus in manu occidentium te 28:10 morte incircumcisorum morieris in manu 28:10 Thou shalt die the deaths of the uncircumcised by

the hand of strangers: for I have spoken it, saith the Lord GOD.

Biblia Sacra Vulgata

28:11 Moreover the word of the LORD came unto me, alienorum quia ego locutus sum ait Dominus Deus 28:11

saying,

et factus est sermo Domini ad me dicens fili hominis leua

28:12 Son of man, take up a lamentation upon the king of planctum super regem Tyri 28:12 et dices ei haec dicit Do-

Tyrus, and say unto him, Thus saith the Lord GOD; Thou sealest up the sum, full of wisdom, and perfect in beauty.

minus Deus tu signaculum similitudinis plenus sapientia et

28:13 Thou hast been in Eden the garden of God; every perfectus decore 28:13 in deliciis paradisi Dei fuisti omnis

precious stone was thy covering, the sardius, topaz, and the diamond, the beryl, the onyx, and the jasper, the sapphire, the emerald, and the carbuncle, and gold: the workmanship of thy tabrets and of thy pipes was prepared in thee in the day that thou wast created.

lapis pretiosus operimentum tuum sardius topazius et iaspis chrysolitus et onyx et berillus sapphyrus et carbunculus et zmaragdus aurum opus decoris tui et foramina tua in die

28:14 Thou art the anointed cherub that covereth; and I qua conditus es praeparata sunt 28:14 tu cherub extentus

have set thee so: thou wast upon the holy mountain of God; thou hast walked up and down in the midst of the

stones of fire. et protegens et posui te in monte sancto Dei in medio lapi-28:15 Thou wast perfect in thy ways from the day that thou dum ignitorum ambulasti 28:15 perfectus in uiis tuis a die

wast created, till iniquity was found in thee.

28:16 By the multitude of thy merchandise they have filled conditionis tuae donec inuenta est iniquitas in te 28:16 in

the midst of thee with violence, and thou hast sinned: therefore I will cast thee as profane out of the mountain of God: and I will destroy thee, O covering cherub, from the midst of the stones of fire.

multitudine negotiationis tuae repleta sunt interiora tua iniquitate et peccasti et eieci te de monte Dei et perdidi te o

28:17 Thine heart was lifted up because of thy beauty, thou cherub protegens de medio lapidum ignitorum 28:17 eleuahast corrupted thy wisdom by reason of thy brightness: I will cast thee to the ground, I will lay thee before kings,

that they may behold thee. tum est cor tuum in decore tuo perdidisti sapientiam tuam

in decore tuo in terram proieci te ante faciem regum dedi te

28:18 Thou hast defiled thy sanctuaries by the multitude ut cernerent te 28:18 in multitudine iniquitatum tuarum et

of thine iniquities, by the iniquity of thy traffick; therefore will I bring forth a fire from the midst of thee, it shall devour thee, and I will bring thee to ashes upon the earth in the sight of all them that behold thee.

iniquitate negotiationis tuae polluisti sanctificationem tuam producam ergo ignem de medio tui qui comedat te et dabo te in cinerem super terram in conspectu omnium uidentium te

28:19 All they that know thee among the people shall be 28:19 omnes qui uiderint te in gentibus obstupescent super te

astonished at thee: thou shalt be a terror, and never shalt thou be any more.

nihili factus es et non eris in perpetuum 28:20 et factus est

28:20 Again the word of the LORD came unto me, saying, 28:21 Son of man, set thy face against Zidon, and proph- sermo Domini ad me dicens 28:21 fili hominis pone faciem esy against it,

28:22 And say, Thus saith the Lord GOD; Behold, I am tuam contra Sidonem et prophetabis de ea 28:22 et dices

against thee, O Zidon; and I will be glorified in the midst of thee: and they shall know that I am the LORD, when I shall have executed judgments in her, and shall be sanctified in her.

haec dicit Dominus Deus ecce ego ad te Sidon et glorificabor in medio tui et scient quia ego Dominus cum fecero in ea

28:23 For I will send into her pestilence, and blood into iudicia et sanctificatus fuero in ea 28:23 et inmittam ei pes-

her streets; and the wounded shall be judged in the midst of her by the sword upon her on every side; and they shall

know that I am the LORD. tilentiam et sanguinem in plateis eius et corruent interfecti in

medio eius gladio per circuitum et scient quia ego Dominus

28:24 And there shall be no more a pricking brier unto the 28:24 et non erit ultra domui Israhel offendiculum amaritudi-

house of Israel, nor any grieving thorn of all that are round about them, that despised them; and they shall know that I

am the Lord GOD. nis et spina dolorem inferens undique per circuitum eorum 28:25 Thus saith the Lord GOD; When I shall have gath- qui aduersantur eis et scient quia ego Dominus Deus 28:25

ered the house of Israel from the people among whom they are scattered, and shall be sanctified in them in the sight of the heathen, then shall they dwell in their land that I have given to my servant Jacob.

Prophetia Ezechielis

haec dicit Dominus Deus quando congregauero domum Israhel de populis in quibus dispersi sunt sanctificabor in eis coram gentibus et habitabunt in terra sua quam dedi seruo

meo Iacob 28:26 et habitabunt in ea securi et aedificabunt 28:26 And they shall dwell safely therein, and shall build houses, and plant vineyards; yea, they shall dwell with confidence, when I have executed judgments upon all those that despise them round about them; and they shall know that I am the LORD their God.

domos plantabuntque uineas et habitabunt confidenter cum fecero iudicia in omnibus qui aduersantur eis per circuitum et scient quia ego Dominus Deus eorum

29:1 in anno decimo in decimo mense undecima mensis 29:1 In the tenth year, in the tenth month, in the twelfth day of the month, the word of the LORD came unto me, saying,

factum est uerbum Domini ad me dicens 29:2 fili hominis

29:2 Son of man, set thy face against Pharaoh king of Egypt, and prophesy against him, and against all Egypt:

pone faciem tuam contra Pharaonem regem Aegypti et prophetabis de eo et de Aegypto uniuersa 29:3 loquere et dices 29:3 Speak, and say, Thus saith the Lord GOD; Behold, I

am against thee, Pharaoh king of Egypt, the great dragon that lieth in the midst of his rivers, which hath said, My

haec dicit Dominus Deus ecce ego ad te Pharao rex Aegypti river is mine own, and I have made it for myself. draco magne qui cubas in medio fluminum tuorum et dicis

meus est fluuius et ego feci memet ipsum 29:4 et ponam 29:4 But I will put hooks in thy jaws, and I will cause the fish of thy rivers to stick unto thy scales, and I will bring thee up out of the midst of thy rivers, and all the fish of thy

frenum in maxillis tuis et adglutinabo pisces fluminum tuo- rivers shall stick unto thy scales. rum squamis tuis et extraham te de medio fluminum tuorum et uniuersi pisces tui squamis tuis adherebunt 29:5 et 29:5 And I will leave thee thrown into the wilderness, thee

and all the fish of thy rivers: thou shalt fall upon the open fields; thou shalt not be brought together, nor gathered: I have given thee for meat to the beasts of the field and to the fowls of the heaven.

proiciam te in desertum et omnes pisces fluminis tui super faciem terrae cades non colligeris neque congregaberis bestiis terrae et uolatilibus caeli dedi te ad deuorandum 29:6 et 29:6 And all the inhabitants of Egypt shall know that I am

the LORD, because they have been a staff of reed to the house of Israel.

scient omnes habitatores Aegypti quia ego Dominus pro eo quod fuisti baculus harundineus domui Israhel 29:7 quando 29:7 When they took hold of thee by thy hand, thou didst

break, and rend all their shoulder: and when they leaned upon thee, thou brakest, and madest all their loins to be at

adprehenderunt te manu et confractus es et lacerasti omnem a stand. umerum eorum et innitentibus eis super te comminutus es et dissoluisti omnes renes eorum 29:8 propterea haec dicit 29:8 Therefore thus saith the Lord GOD; Behold, I will

bring a sword upon thee, and cut off man and beast out of thee.

Dominus Deus ecce ego adducam super te gladium et inter-

ficiam de te hominem et iumentum 29:9 et erit terra Ae- 29:9 And the land of Egypt shall be desolate and waste; and they shall know that I am the LORD: because he hath said, The river is mine, and I have made it.

gypti in desertum et solitudinem et scient quia ego Dominus

eo quod dixerit fluuius meus est et ego feci 29:10 idcirco 29:10 Behold, therefore I am against thee, and against thy rivers, and I will make the land of Egypt utterly waste and desolate, from the tower of Syene even unto the border of

ecce ego ad te et ad flumina tua daboque terram Aegypti in Ethiopia. solitudines gladio dissipatam a turre Syenes usque ad terminos Aethiopiae 29:11 non pertransibit eam pes hominis 29:11 No foot of man shall pass through it, nor foot of

beast shall pass through it, neither shall it be inhabited forty years.

neque pes iumenti gradietur in ea et non habitabitur quadra-

Biblia Sacra Vulgata

29:12 And I will make the land of Egypt desolate in the ginta annis 29:12 daboque terram Aegypti desertam in me-

midst of the countries that are desolate, and her cities among the cities that are laid waste shall be desolate forty years: and I will scatter the Egyptians among the nations, and will disperse them through the countries.

dio terrarum desertarum et ciuitates eius in medio urbium subuersarum erunt desolatae quadraginta annis et dispergam

29:13 Yet thus saith the Lord GOD; At the end of forty Aegyptios in nationes et uentilabo eos in terras 29:13 quia

years will I gather the Egyptians from the people whither they were scattered:

haec dicit Dominus Deus post finem quadraginta annorum congregabo Aegyptum de populis in quibus dispersi fuerunt

29:14 And I will bring again the captivity of Egypt, and 29:14 et reducam captiuitatem Aegypti et conlocabo eos in

will cause them to return into the land of Pathros, into the land of their habitation; and they shall be there a base king-

dom. terra Fatures in terra natiuitatis suae et erunt ibi in regnum

29:15 It shall be the basest of the kingdoms; neither shall it humile 29:15 inter regna cetera erit humillima et non eleexalt itself any more above the nations: for I will diminish them, that they shall no more rule over the nations.

uabitur ultra super nationes et inminuam eos ne imperent

29:16 And it shall be no more the confidence of the house gentibus 29:16 neque erunt ultra domui Israhel in confiden-

of Israel, which bringeth their iniquity to remembrance, when they shall look after them: but they shall know that

I am the Lord GOD. tia docentes iniquitatem ut fugiant et sequantur eos et scient

29:17 And it came to pass in the seven and twentieth year, quia ego Dominus Deus 29:17 et factum est in uicesimo et in the first month, in the first day of the month, the word of the LORD came unto me, saying,

septimo anno in primo in una mensis factum est uerbum Do-

29:18 Son of man, Nebuchadrezzar king of Babylon mini ad me dicens 29:18 fili hominis Nabuchodonosor rex

caused his army to serve a great service against Tyrus: every head was made bald, and every shoulder was peeled: yet had he no wages, nor his army, for Tyrus, for the service that he had served against it:

Babylonis seruire fecit exercitum suum seruitute magna aduersum Tyrum omne caput decaluatum et omnis umerus depilatus est et merces non est reddita ei neque exercitui eius

29:19 Therefore thus saith the Lord GOD; Behold, I will de Tyro pro seruitute qua seruiuit mihi aduersum eam 29:19

give the land of Egypt unto Nebuchadrezzar king of Babylon; and he shall take her multitude, and take her spoil, and

take her prey; and it shall be the wages for his army. propterea haec dicit Dominus Deus ecce ego dabo Nabuchodonosor regem Babylonis in terra Aegypti et accipiet multitudinem eius et depraedabitur manubias eius et diripiet spo-

29:20 I have given him the land of Egypt for his labour lia eius et erit merces exercitui illius 29:20 et operi pro quo

wherewith he served against it, because they wrought for me, saith the Lord GOD.

seruiuit aduersum eam dedi ei terram Aegypti pro eo quod

29:21 In that day will I cause the horn of the house of laborauerunt mihi ait Dominus Deus 29:21 in die illo pul-

Israel to bud forth, and I will give thee the opening of the mouth in the midst of them; and they shall know that I am

the LORD. lulabit cornu domui Israhel et tibi dabo apertum os in medio eorum et scient quoniam ego Dominus

30:1 The word of the LORD came again unto me, saying, 30:1 et factum est uerbum Domini ad me dicens 30:2 30:2 Son of man, prophesy and say, Thus saith the Lord GOD; Howl ye, Woe worth the day!

fili hominis propheta et dic haec dicit Dominus Deus ulu-

30:3 For the day is near, even the day of the LORD is near, late uae uae diei 30:3 quia iuxta est dies et adpropinquauit

a cloudy day; it shall be the time of the heathen.

30:4 And the sword shall come upon Egypt, and great pain dies Domini dies nubis tempus gentium erit 30:4 et ueniet

shall be in Ethiopia, when the slain shall fall in Egypt, and they shall take away her multitude, and her foundations

shall be broken down. gladius in Aegyptum et erit pauor in Aethiopia cum ceciderint uulnerati in Aegypto et ablata fuerit multitudo illius et

Prophetia Ezechielis

destructa fundamenta eius 30:5 Aethiopia et Lybia et Ly- 30:5 Ethiopia, and Libya, and Lydia, and all the mingled

people, and Chub, and the men of the land that is in league, shall fall with them by the sword.

dii et omne reliquum uulgus et Chub et filii terrae foederis cum eis gladio cadent 30:6 haec dicit Dominus Deus et 30:6 Thus saith the LORD; They also that uphold Egypt

shall fall; and the pride of her power shall come down: from the tower of Syene shall they fall in it by the sword,

corruent fulcientes Aegyptum et destruetur superbia imperii saith the Lord GOD. eius a turre Syenes gladio cadent in ea ait Dominus exercituum 30:7 et dissipabuntur in medio terrarum desolatarum 30:7 And they shall be desolate in the midst of the coun-

tries that are desolate, and her cities shall be in the midst of the cities that are wasted.

et urbes eius in medio ciuitatum desertarum erunt 30:8 et

30:8 And they shall know that I am the LORD, when I have set a fire in Egypt, and when all her helpers shall be destroyed.

scient quoniam ego Dominus cum dedero ignem in Aegyptum et adtriti fuerint omnes auxiliatores eius 30:9 in die 30:9 In that day shall messengers go forth from me in ships

to make the careless Ethiopians afraid, and great pain shall come upon them, as in the day of Egypt: for, lo, it cometh.

illa egredientur nuntii a facie mea in trieribus ad conterendam Aethiopiae confidentiam et erit pauor in eis in die Aegypti quia absque dubio ueniet 30:10 haec dicit Dominus 30:10 Thus saith the Lord GOD; I will also make the mul-

titude of Egypt to cease by the hand of Nebuchadrezzar king of Babylon.

Deus et cessare faciam multitudinem Aegypti in manu Nabuchodonosor regis Babylonis 30:11 ipse et populus eius 30:11 He and his people with him, the terrible of the na-

tions, shall be brought to destroy the land: and they shall draw their swords against Egypt, and fill the land with the

cum eo fortissimi gentium adducentur ad disperdendam ter- slain. ram et euaginabunt gladios suos super Aegyptum et implebunt terram interfectis 30:12 et faciam alueos fluminum 30:12 And I will make the rivers dry, and sell the land

into the hand of the wicked: and I will make the land waste, and all that is therein, by the hand of strangers: I

aridos et tradam terram in manu pessimorum et dissipabo the LORD have spoken it. terram et plenitudinem eius in manu alienorum ego Dominus locutus sum 30:13 haec dicit Dominus Deus et disper- 30:13 Thus saith the Lord GOD; I will also destroy the

idols, and I will cause their images to cease out of Noph; and there shall be no more a prince of the land of Egypt:

dam simulacra et cessare faciam idola de Memphis et dux and I will put a fear in the land of Egypt. de terra Aegypti non erit amplius et dabo terrorem in terra Aegypti 30:14 et disperdam terram Fatures et dabo ignem 30:14 And I will make Pathros desolate, and will set fire

in Zoan, and will execute judgments in No.

in Tafnis et faciam iudicia in Alexandriam 30:15 et effun- 30:15 And I will pour my fury upon Sin, the strength of

Egypt; and I will cut off the multitude of No.

dam indignationem meam super Pelusium robur Aegypti et interficiam multitudinem Alexandriae 30:16 et dabo ignem 30:16 And I will set fire in Egypt: Sin shall have great

pain, and No shall be rent asunder, and Noph shall have distresses daily.

in Aegypto quasi parturiens dolebit Pelusium et Alexandria erit dissipata et in Memphis angustiae cotidianae 30:17 iu- 30:17 The young men of Aven and of Pibeseth shall fall

by the sword: and these cities shall go into captivity.

uenes Eliupoleos et Bubasti gladio cadent et ipsae captiuae ducentur 30:18 et in Tafnis nigrescet dies cum contriuero 30:18 At Tehaphnehes also the day shall be darkened,

when I shall break there the yokes of Egypt: and the pomp of her strength shall cease in her: as for her, a cloud shall

ibi sceptra Aegypti et defecerit in ea superbia potentiae eius cover her, and her daughters shall go into captivity. ipsam nubes operiet filiae autem eius in captiuitatem ducentur 30:19 et faciam iudicia in Aegypto et scient quia ego 30:19 Thus will I execute judgments in Egypt: and they

shall know that I am the LORD.

Biblia Sacra Vulgata

30:20 And it came to pass in the eleventh year, in the first Dominus 30:20 et factum est in undecimo anno in primo month, in the seventh day of the month, that the word of the LORD came unto me, saying,

in septima mensis factum est uerbum Domini ad me dicens

30:21 Son of man, I have broken the arm of Pharaoh king 30:21 fili hominis brachium Pharao regis Aegypti confregi

of Egypt; and, lo, it shall not be bound up to be healed, to put a roller to bind it, to make it strong to hold the sword.

et ecce non est obuolutum ut restitueretur ei sanitas ut ligaretur pannis et farciretur linteolis et recepto robore posset

30:22 Therefore thus saith the Lord GOD; Behold, I am tenere gladium 30:22 propterea haec dicit Dominus Deus

against Pharaoh king of Egypt, and will break his arms, the strong, and that which was broken; and I will cause

the sword to fall out of his hand. ecce ego ad Pharao regem Aegypti et comminuam brachium

eius forte sed confractum et deiciam gladium de manu eius

30:23 And I will scatter the Egyptians among the nations, 30:23 et dispergam Aegyptum in gentibus et uentilabo eos in

and will disperse them through the countries.

30:24 And I will strengthen the arms of the king of Baby- terris 30:24 et confortabo brachia regis Babylonis daboque

lon, and put my sword in his hand: but I will break Pharaoh’s arms, and he shall groan before him with the

groanings of a deadly wounded man. gladium meum in manu eius et confringam brachia Phara-30:25 But I will strengthen the arms of the king of Baby- onis et gement gemitibus interfecti coram facie eius 30:25

lon, and the arms of Pharaoh shall fall down; and they shall know that I am the LORD, when I shall put my sword into the hand of the king of Babylon, and he shall stretch it out upon the land of Egypt.

et confortabo brachia regis Babylonis et brachia Pharaonis concident et scient quia ego Dominus cum dedero gladium meum in manu regis Babylonis et extenderit eum super ter-

30:26 And I will scatter the Egyptians among the nations, ram Aegypti 30:26 et dispergam Aegyptum in nationes et and disperse them among the countries; and they shall know that I am the LORD.

uentilabo eos in terris et scient quia ego Dominus

31:1 And it came to pass in the eleventh year, in the third 31:1 et factum est in undecimo anno tertio una mensis fac-

month, in the first day of the month, that the word of the LORD came unto me, saying,

tum est uerbum Domini ad me dicens 31:2 fili hominis dic

31:2 Son of man, speak unto Pharaoh king of Egypt, and to his multitude; Whom art thou like in thy greatness?

Pharaoni regi Aegypti et populo eius cui similis factus es

31:3 Behold, the Assyrian was a cedar in Lebanon with in magnitudine tua 31:3 ecce Assur quasi cedrus in Libano

fair branches, and with a shadowing shroud, and of an high stature; and his top was among the thick boughs.

pulcher ramis et frondibus nemorosus excelsusque altitudine

31:4 The waters made him great, the deep set him up on et inter condensas frondes eleuatum est cacumen eius 31:4

high with her rivers running round about his plants, and sent her little rivers unto all the trees of the field.

aquae nutrierunt illum abyssus exaltauit eum flumina eius manabant in circuitu radicum eius et riuos suos emisit ad

31:5 Therefore his height was exalted above all the trees of uniuersa ligna regionis 31:5 propterea eleuata est altitudo

the field, and his boughs were multiplied, and his branches became long because of the multitude of waters, when he

shot forth. eius super omnia ligna regionis et multiplicata sunt arbusta 31:6 All the fowls of heaven made their nests in his eius et eleuati sunt rami eius prae aquis multis 31:6 cumque

boughs, and under his branches did all the beasts of the field bring forth their young, and under his shadow dwelt

all great nations. extendisset umbram suam in ramis eius fecerunt nidos omnia uolatilia caeli et sub frondibus eius genuerunt omnes bestiae saltuum et sub umbraculo illius habitabat coetus gentium

31:7 Thus was he fair in his greatness, in the length of his plurimarum 31:7 eratque pulcherrimus in magnitudine sua

branches: for his root was by great waters.

et in dilatatione arbustorum suorum erat enim radix illius

Prophetia Ezechielis

iuxta aquas multas 31:8 cedri non fuerunt altiores illo in 31:8 The cedars in the garden of God could not hide him:

the fir trees were not like his boughs, and the chestnut trees were not like his branches; nor any tree in the garden of

paradiso Dei abietes non adaequauerunt summitatem eius et God was like unto him in his beauty. platani non fuerunt aequae frondibus illius omne lignum paradisi Dei non est adsimilatum illi et pulchritudini eius 31:9 31:9 I have made him fair by the multitude of his branches:

so that all the trees of Eden, that were in the garden of God, envied him.

quoniam speciosum feci eum et multis condensisque frondibus et aemulata sunt eum omnia ligna uoluptatis quae erant in paradiso Dei 31:10 propterea haec dicit Dominus Deus 31:10 Therefore thus saith the Lord GOD; Because thou

hast lifted up thyself in height, and he hath shot up his top among the thick boughs, and his heart is lifted up in his

pro eo quod sublimatus est in altitudine et dedit summita- height; tem suam uirentem atque condensam et eleuatum est cor eius in altitudine sua 31:11 tradidi eum in manu fortissimi gen- 31:11 I have therefore delivered him into the hand of the

mighty one of the heathen; he shall surely deal with him: I have driven him out for his wickedness.

tium faciens faciet ei iuxta impietatem eius eieci eum 31:12

31:12 And strangers, the terrible of the nations, have cut him off, and have left him: upon the mountains and in all the valleys his branches are fallen, and his boughs are broken by all the rivers of the land; and all the people of the earth are gone down from his shadow, and have left him.

et succident illum alieni et crudelissimi nationum et proicient eum super montes et in cunctis conuallibus corruent rami eius et confringentur arbusta eius in uniuersis rupibus terrae et recedent de umbraculo eius omnes populi terrae et relinquent eum 31:13 in ruina eius habitauerunt omnia uo- 31:13 Upon his ruin shall all the fowls of the heaven re-

main, and all the beasts of the field shall be upon his branches:

latilia caeli et in ramis eius fuerunt uniuersae bestiae regionis 31:14 quam ob rem non eleuabuntur in altitudine sua 31:14 To the end that none of all the trees by the waters

exalt themselves for their height, neither shoot up their top among the thick boughs, neither their trees stand up in their height, all that drink water: for they are all delivered unto death, to the nether parts of the earth, in the midst of the children of men, with them that go down to the pit.

omnia ligna aquarum neque ponent sublimitatem suam inter nemorosa atque frondosa nec stabunt in sublimitate eorum omnia quae inrigantur aquis quia omnes traditi sunt in mortem ad terram ultimam in medio filiorum hominum ad eos qui descendunt in lacum 31:15 haec dicit Dominus De- 31:15 Thus saith the Lord GOD; In the day when he went

down to the grave I caused a mourning: I covered the deep for him, and I restrained the floods thereof, and the great waters were stayed: and I caused Lebanon to mourn for him, and all the trees of the field fainted for him.

us in die quando descendit ad inferos indixi luctum operui eum abysso et prohibui flumina eius et coercui aquas multas contristatus est super eum Libanus et omnia ligna agri concussa sunt 31:16 a sonitu ruinae eius commoui gentes cum 31:16 I made the nations to shake at the sound of his fall,

when I cast him down to hell with them that descend into the pit: and all the trees of Eden, the choice and best of Lebanon, all that drink water, shall be comforted in the nether parts of the earth.

deducerem eum ad infernum cum his qui descendebant in lacum et consolata sunt in terra infima omnia ligna uoluptatis egregia atque praeclara in Libano uniuersa quae inrigabantur aquis 31:17 nam et ipsi cum ea descendent ad infernum 31:17 They also went down into hell with him unto them

that be slain with the sword; and they that were his arm, that dwelt under his shadow in the midst of the heathen.

ad interfectos gladio et brachium uniuscuiusque sedebit sub umbraculo eius in medio nationum 31:18 cui adsimilatus es 31:18 To whom art thou thus like in glory and in greatness

among the trees of Eden? yet shalt thou be brought down with the trees of Eden unto the nether parts of the earth: thou shalt lie in the midst of the uncircumcised with them that be slain by the sword. This is Pharaoh and all his multitude, saith the Lord GOD.

Biblia Sacra Vulgata

o inclite atque sublimis inter ligna uoluptatis ecce deductus es cum lignis uoluptatis ad terram ultimam in medio incircumcisorum dormies cum his qui interfecti sunt gladio ipse est Pharao et omnis multitudo eius dicit Dominus Deus

32:1 And it came to pass in the twelfth year, in the twelfth 32:1 et factum est duodecimo anno in mense duodecimo in month, in the first day of the month, that the word of the LORD came unto me, saying,

una mensis factum est uerbum Domini ad me dicens 32:2

32:2 Son of man, take up a lamentation for Pharaoh king of Egypt, and say unto him, Thou art like a young lion of the nations, and thou art as a whale in the seas: and thou camest forth with thy rivers, and troubledst the waters with thy feet, and fouledst their rivers.

fili hominis adsume lamentum super Pharao regem Aegypti et dices ad eum leoni gentium adsimilatus es et draconi qui est in mari et uentilabas cornu in fluminibus tuis et conturba-

32:3 Thus saith the Lord GOD; I will therefore spread out bas aquas pedibus tuis et conculcabas flumina eorum 32:3 my net over thee with a company of many people; and they shall bring thee up in my net.

propterea haec dicit Dominus Deus expandam super te rete meum in multitudine populorum multorum et extrahent te in

32:4 Then will I leave thee upon the land, I will cast thee sagena mea 32:4 et proiciam te in terram super faciem agri

forth upon the open field, and will cause all the fowls of the heaven to remain upon thee, and I will fill the beasts of

the whole earth with thee. abiciam te et habitare faciam super te omnia uolatilia caeli 32:5 And I will lay thy flesh upon the mountains, and fill et saturabo de te bestias uniuersae terrae 32:5 et dabo car-

the valleys with thy height.

32:6 I will also water with thy blood the land wherein thou nes tuas super montes et implebo colles tuos sanie tua 32:6

swimmest, even to the mountains; and the rivers shall be full of thee.

et inrigabo terram pedore sanguinis tui super montes et ual-

32:7 And when I shall put thee out, I will cover the heaven, les implebuntur ex te 32:7 et operiam cum extinctus fueris

and make the stars thereof dark; I will cover the sun with a cloud, and the moon shall not give her light.

caelos et nigrescere faciam stellas eius solem nube tegam

32:8 All the bright lights of heaven will I make dark over et luna non dabit lumen suum 32:8 omnia luminaria caeli thee, and set darkness upon thy land, saith the Lord GOD.

maerere faciam super te et dabo tenebras super terram tuam

32:9 I will also vex the hearts of many people, when I shall dicit Dominus Deus 32:9 et inritabo cor populorum multobring thy destruction among the nations, into the countries which thou hast not known.

rum cum induxero contritionem tuam in gentibus super ter-

32:10 Yea, I will make many people amazed at thee, and ras quas nescis 32:10 et stupescere faciam super te populos

their kings shall be horribly afraid for thee, when I shall brandish my sword before them; and they shall tremble at every moment, every man for his own life, in the day of thy fall.

multos et reges eorum horrore nimio formidabunt super te cum uolare coeperit gladius meus super facies eorum et obstupescent repente singuli pro anima sua in die ruinae suae

32:11 For thus saith the Lord GOD; The sword of the king 32:11 quia haec dicit Dominus Deus gladius regis Babylo-

of Babylon shall come upon thee.

32:12 By the swords of the mighty will I cause thy mul- nis ueniet tibi 32:12 in gladiis fortium deiciam multitudi-

titude to fall, the terrible of the nations, all of them: and they shall spoil the pomp of Egypt, and all the multitude

thereof shall be destroyed. nem tuam inexpugnabiles gentes omnes heae et uastabunt

32:13 I will destroy also all the beasts thereof from beside superbiam Aegypti et dissipabitur multitudo eius 32:13 et the great waters; neither shall the foot of man trouble them any more, nor the hoofs of beasts trouble them.

perdam omnia iumenta eius quae erant super aquas plurimas et non conturbabit eas pes hominis ultra neque ungula iu-

Prophetia Ezechielis

mentorum turbabit eas 32:14 tunc purissimas reddam aquas 32:14 Then will I make their waters deep, and cause their

rivers to run like oil, saith the Lord GOD.

eorum et flumina eorum quasi oleum adducam ait Dominus Deus 32:15 cum dedero terram Aegypti desolatam desere- 32:15 When I shall make the land of Egypt desolate, and

the country shall be destitute of that whereof it was full, when I shall smite all them that dwell therein, then shall

tur autem terra a plenitudine sua quando percussero omnes they know that I am the LORD.

habitatores eius et scient quia ego Dominus 32:16 planc- 32:16 This is the lamentation wherewith they shall lament her: the daughters of the nations shall lament her: they shall lament for her, even for Egypt, and for all her multi-

tus est et plangent eum filiae gentium plangent eum super tude, saith the Lord GOD. Aegypto et super multitudine eius plangent eum ait Domi-

nus Deus 32:17 et factum est in duodecimo anno in quin- 32:17 It came to pass also in the twelfth year, in the fifteenth day of the month, that the word of the LORD came unto me, saying,

tadecima mensis factum est uerbum Domini ad me dicens 32:18 fili hominis cane lugubre super multitudine Aegypti et 32:18 Son of man, wail for the multitude of Egypt, and

cast them down, even her, and the daughters of the famous nations, unto the nether parts of the earth, with them that

detrahe eam ipsam et filias gentium robustarum ad terram go down into the pit. ultimam cum his qui descendunt in lacum 32:19 quo pulch- 32:19 Whom dost thou pass in beauty? go down, and be

thou laid with the uncircumcised.

rior es descende et dormi cum incircumcisis 32:20 in me- 32:20 They shall fall in the midst of them that are slain by the sword: she is delivered to the sword: draw her and all her multitudes.

dio interfectorum gladio cadent gladius datus est adtraxerunt eam et omnes populos eius 32:21 loquentur ei poten- 32:21 The strong among the mighty shall speak to him out

of the midst of hell with them that help him: they are gone down, they lie uncircumcised, slain by the sword.

tissimi robustorum de medio inferni qui cum auxiliatoribus eius descenderunt et dormierunt incircumcisi interfecti gladio 32:22 ibi Assur et omnis multitudo eius in circuitu il- 32:22 Asshur is there and all her company: his graves are

about him: all of them slain, fallen by the sword:

lius sepulchra eius omnes interfecti et qui ceciderunt gladio 32:23 quorum data sunt sepulchra in nouissimis laci et facta 32:23 Whose graves are set in the sides of the pit, and her

company is round about her grave: all of them slain, fallen by the sword, which caused terror in the land of the living.

est multitudo eius per gyrum sepulchri eius uniuersi interfecti cadentesque gladio qui dederant quondam formidinem in terra uiuentium 32:24 ibi Aelam et omnis multitudo eius 32:24 There is Elam and all her multitude round about

her grave, all of them slain, fallen by the sword, which are gone down uncircumcised into the nether parts of the earth, which caused their terror in the land of the living; yet have they borne their shame with them that go down to the pit.

per gyrum sepulchri sui omnes hii interfecti ruentesque gladio qui descenderunt incircumcisi ad terram ultimam qui posuerunt terrorem suum in terra uiuentium et portauerunt ignominiam suam cum his qui descendunt in lacum 32:25 in 32:25 They have set her a bed in the midst of the slain

with all her multitude: her graves are round about him: all of them uncircumcised, slain by the sword: though their terror was caused in the land of the living, yet have they borne their shame with them that go down to the pit: he is put in the midst of them that be slain.

medio interfectorum posuerunt cubile eius in uniuersis populis eius in circuitu eius sepulchrum illius omnes hii incircumcisi interfectique gladio dederant enim terrorem in terra uiuentium et portauerunt ignominiam suam cum his qui descendunt in lacum in medio interfectorum positi sunt 32:26 32:26 There is Meshech, Tubal, and all her multitude: her

graves are round about him: all of them uncircumcised, slain by the sword, though they caused their terror in the

ibi Mosoch et Thubal et omnis multitudo eius in circuitu il- land of the living.

Biblia Sacra Vulgata

lius sepulchra eius omnes hii incircumcisi interfectique et cadentes gladio quia dederunt formidinem suam in terra ui-

32:27 And they shall not lie with the mighty that are fallen uentium 32:27 et non dormient cum fortibus cadentibusque of the uncircumcised, which are gone down to hell with their weapons of war: and they have laid their swords under their heads, but their iniquities shall be upon their bones, though they were the terror of the mighty in the land of the living.

et incircumcisis qui descenderunt ad infernum cum armis suis et posuerunt gladios suos sub capitibus suis et fuerunt iniquitates eorum in ossibus eorum quia terror fortium facti

32:28 Yea, thou shalt be broken in the midst of the uncir- sunt in terra uiuentium 32:28 et tu ergo in medio incircum-

cumcised, and shalt lie with them that are slain with the sword.

cisorum contereris et dormies cum interfectis gladio 32:29

32:29 There is Edom, her kings, and all her princes, which with their might are laid by them that were slain by the sword: they shall lie with the uncircumcised, and with them that go down to the pit.

ibi Idumea et reges eius omnes duces eius qui dati sunt cum exercitu suo cum interfectis gladio et qui cum incircumcisis

32:30 There be the princes of the north, all of them, and all dormierunt et cum his qui descenderunt in lacum 32:30 ibi the Zidonians, which are gone down with the slain; with their terror they are ashamed of their might; and they lie uncircumcised with them that be slain by the sword, and bear their shame with them that go down to the pit.

principes aquilonis omnes et uniuersi uenatores qui deducti sunt cum interfectis pauentes et in sua fortitudine confusi qui dormierunt incircumcisi cum interfectis gladio et portauerunt confusionem suam cum his qui descendunt in la-

32:31 Pharaoh shall see them, and shall be comforted over cum 32:31 uidit eos Pharao et consolatus est super uniuersa

all his multitude, even Pharaoh and all his army slain by the sword, saith the Lord GOD.

multitudine sua quae interfecta est gladio Pharao et omnis

32:32 For I have caused my terror in the land of the living: exercitus eius ait Dominus Deus 32:32 quia dedi terrorem and he shall be laid in the midst of the uncircumcised with them that are slain with the sword, even Pharaoh and all

his multitude, saith the Lord GOD. meum in terra uiuentium et dormiuit in medio incircumcisorum cum interfectis gladio Pharao et omnis multitudo eius ait Dominus Deus

33:1 Again the word of the LORD came unto me, saying, 33:1 et factum est uerbum Domini ad me dicens 33:2 fili

33:2 Son of man, speak to the children of thy people, and say unto them, When I bring the sword upon a land, if the people of the land take a man of their coasts, and set him for their watchman:

hominis loquere ad filios populi tui et dices ad eos terra cum induxero super eam gladium et tulerit populus terrae uirum unum de nouissimis suis et constituerit eum super se specu-

33:3 If when he seeth the sword come upon the land, he latorem 33:3 et ille uiderit gladium uenientem super terram

blow the trumpet, and warn the people;

33:4 Then whosoever heareth the sound of the trumpet, et cecinerit bucina et adnuntiauerit populo 33:4 audiens au-

and taketh not warning; if the sword come, and take him away, his blood shall be upon his own head.

tem quisquis ille est sonum bucinae non se obseruauerit ueneritque gladius et tulerit eum sanguis ipsius super caput

33:5 He heard the sound of the trumpet, and took not warn- eius erit 33:5 sonum bucinae audiuit et non se obseruauit

ing; his blood shall be upon him. But he that taketh warning shall deliver his soul.

sanguis eius in ipso erit si autem custodierit animam suam

33:6 But if the watchman see the sword come, and blow saluauit 33:6 quod si speculator uiderit gladium uenientem

not the trumpet, and the people be not warned; if the sword come, and take any person from among them, he is taken away in his iniquity; but his blood will I require at the watchman’s hand.

et non insonuerit bucina et populus non se custodierit uene-

Prophetia Ezechielis

ritque gladius et tulerit de eis animam ille quidem in iniquitate sua captus est sanguinem autem eius de manu speculatoris requiram 33:7 et tu fili hominis speculatorem dedi te 33:7 So thou, O son of man, I have set thee a watchman

unto the house of Israel; therefore thou shalt hear the word at my mouth, and warn them from me.

domui Israhel audiens ergo ex ore meo sermonem adnuntiabis eis ex me 33:8 si me dicente ad impium impie morte 33:8 When I say unto the wicked, O wicked man, thou

shalt surely die; if thou dost not speak to warn the wicked from his way, that wicked man shall die in his iniquity; but

morieris non fueris locutus ut se custodiat impius a uia sua his blood will I require at thine hand. ipse impius in iniquitate sua morietur sanguinem autem eius de manu tua requiram 33:9 si autem adnuntiante te ad im- 33:9 Nevertheless, if thou warn the wicked of his way to

turn from it; if he do not turn from his way, he shall die in his iniquity; but thou hast delivered thy soul.

pium ut a uiis suis conuertatur non fuerit conuersus a uia sua ipse in iniquitate sua morietur porro tu animam tuam liberasti 33:10 tu ergo fili hominis dic ad domum Israhel sic lo- 33:10 Therefore, O thou son of man, speak unto the house

of Israel; Thus ye speak, saying, If our transgressions and our sins be upon us, and we pine away in them, how should

cuti estis dicentes iniquitates nostrae et peccata nostra super we then live? nos sunt et in ipsis nos tabescimus quomodo ergo uiuere poterimus 33:11 dic ad eos uiuo ego dicit Dominus Deus nolo 33:11 Say unto them, As I live, saith the Lord GOD, I have

no pleasure in the death of the wicked; but that the wicked turn from his way and live: turn ye, turn ye from your evil

mortem impii sed ut reuertatur impius a uia sua et uiuat con- ways; for why will ye die, O house of Israel? uertimini a uiis uestris pessimis et quare moriemini domus Israhel 33:12 tu itaque fili hominis dic ad filios populi tui 33:12 Therefore, thou son of man, say unto the children

of thy people, The righteousness of the righteous shall not deliver him in the day of his transgression: as for the wickedness of the wicked, he shall not fall thereby in the day that he turneth from his wickedness; neither shall the righteous be able to live for his righteousness in the day that he sinneth.

iustitia iusti non liberabit eum in quacumque die peccauerit et impietas impii non nocebit ei in quacumque die conuersus fuerit ab impietate sua et iustus non poterit uiuere in iustitia sua in quacumque die peccauerit 33:13 etiam si dixero iusto 33:13 When I shall say to the righteous, that he shall surely

live; if he trust to his own righteousness, and commit iniquity, all his righteousnesses shall not be remembered; but

quod uita uiuat et confisus in iustitia sua fecerit iniquitatem for his iniquity that he hath committed, he shall die for it. omnes iustitiae eius obliuioni tradentur et in iniquitate sua quam operatus est in ipsa morietur 33:14 sin autem dixero 33:14 Again, when I say unto the wicked, Thou shalt

surely die; if he turn from his sin, and do that which is lawful and right;

impio morte morieris et egerit paenitentiam a peccato suo feceritque iudicium et iustitiam 33:15 pignus restituerit ille 33:15 If the wicked restore the pledge, give again that he

had robbed, walk in the statutes of life, without committing iniquity; he shall surely live, he shall not die.

impius rapinamque reddiderit in mandatis uitae ambulauerit nec fecerit quicquam iniustum uita uiuet et non morietur 33:16 omnia peccata eius quae peccauit non inputabuntur ei 33:16 None of his sins that he hath committed shall be

mentioned unto him: he hath done that which is lawful and right; he shall surely live.

iudicium et iustitiam fecit uita uiuet 33:17 et dixerunt filii

33:17 Yet the children of thy people say, The way of the Lord is not equal: but as for them, their way is not equal.

populi tui non est aequi ponderis uia Domini et ipsorum uia iniusta est 33:18 cum enim recesserit iustus a iustitia sua 33:18 When the righteous turneth from his righteousness,

and committeth iniquity, he shall even die thereby.

feceritque iniquitatem morietur in eis 33:19 et cum recesse- 33:19 But if the wicked turn from his wickedness, and do

that which is lawful and right, he shall live thereby.

Biblia Sacra Vulgata

rit impius ab impietate sua feceritque iudicium et iustitiam

33:20 Yet ye say, The way of the Lord is not equal. O ye uiuet in eis 33:20 et dicitis non est recta uia Domini unum-

house of Israel, I will judge you every one after his ways.

quemque iuxta uias suas iudicabo de uobis domus Israhel

33:21 And it came to pass in the twelfth year of our cap- 33:21 et factum est in duodecimo anno in duodecimo mense

tivity, in the tenth month, in the fifth day of the month, that one that had escaped out of Jerusalem came unto me,

saying, The city is smitten. in quinta mensis transmigrationis nostrae uenit ad me qui fu-33:22 Now the hand of the LORD was upon me in the gerat de Hierusalem dicens uastata est ciuitas 33:22 manus

evening, afore he that was escaped came; and had opened my mouth, until he came to me in the morning; and my

mouth was opened, and I was no more dumb. autem Domini facta fuerat ad me uespere antequam ueniret

qui fugerat aperuitque os meum donec ueniret ad me mane

33:23 Then the word of the LORD came unto me, saying, et aperto ore meo non silui amplius 33:23 et factum est uer-

33:24 Son of man, they that inhabit those wastes of the bum Domini ad me dicens 33:24 fili hominis qui habitant land of Israel speak, saying, Abraham was one, and he inherited the land: but we are many; the land is given us

for inheritance. in ruinosis his super humum Israhel loquentes aiunt unus

erat Abraham et hereditate possedit terram nos autem multi

33:25 Wherefore say unto them, Thus saith the Lord nobis data est terra in possessionem 33:25 idcirco dices ad

GOD; Ye eat with the blood, and lift up your eyes toward your idols, and shed blood: and shall ye possess the land?

eos haec dicit Dominus Deus qui in sanguine comeditis et oculos uestros leuatis ad inmunditias uestras et sanguinem

33:26 Ye stand upon your sword, ye work abomination, funditis numquid terram hereditate possidebitis 33:26 steti-

and ye defile every one his neighbour’s wife: and shall ye possess the land?

stis in gladiis uestris fecistis abominationes et unusquisque uxorem proximi sui polluit et terram hereditate possidebi-

33:27 Say thou thus unto them, Thus saith the Lord GOD; tis 33:27 haec dices ad eos sic dicit Dominus Deus uiuo As I live, surely they that are in the wastes shall fall by the sword, and him that is in the open field will I give to the beasts to be devoured, and they that be in the forts and in the caves shall die of the pestilence.

ego quia qui in ruinosis habitant gladio cadent et qui in agro est bestiis tradetur ad deuorandum qui autem in praesidiis et

33:28 For I will lay the land most desolate, and the pomp in speluncis sunt peste morientur 33:28 et dabo terram in

of her strength shall cease; and the mountains of Israel shall be desolate, that none shall pass through.

solitudinem et desertum et deficiet superba fortitudo eius et desolabuntur montes Israhel eo quod nullus sit qui per eos

33:29 Then shall they know that I am the LORD, when I transeat 33:29 et scient quia ego Dominus cum dedero ter-

have laid the land most desolate because of all their abominations which they have committed.

ram desolatam et desertam propter uniuersas abominationes

33:30 Also, thou son of man, the children of thy people suas quas operati sunt 33:30 et tu fili hominis filii populi

still are talking against thee by the walls and in the doors of the houses, and speak one to another, every one to his brother, saying, Come, I pray you, and hear what is the word that cometh forth from the LORD.

tui qui loquuntur de te iuxta muros et in ostiis domorum et dicunt unus ad alterum uir ad proximum suum loquentes ue-

33:31 And they come unto thee as the people cometh, and nite et audiamus qui sit sermo egrediens a Domino 33:31

they sit before thee as my people, and they hear thy words, but they will not do them: for with their mouth they shew much love, but their heart goeth after their covetousness. et ueniunt ad te quasi si ingrediatur populus et sedent co-

ram te populus meus et audiunt sermones tuos et non faciunt eos quia in canticum oris sui uertunt illos et auaritiam suam

Prophetia Ezechielis

sequitur cor eorum 33:32 et es eis quasi carmen musicum 33:32 And, lo, thou art unto them as a very lovely song

of one that hath a pleasant voice, and can play well on an instrument: for they hear thy words, but they do them not.

quod suaui dulcique sono canitur et audient uerba tua et non facient ea 33:33 et cum uenerit quod praedictum est ecce 33:33 And when this cometh to pass, (lo, it will come,)

then shall they know that a prophet hath been among them.

enim uenit tunc scient quod prophetes fuerit inter eos

34:1 et factum est uerbum Domini ad me dicens 34:2 fili 34:1 And the word of the LORD came unto me, saying,

34:2 Son of man, prophesy against the shepherds of Israel, prophesy, and say unto them, Thus saith the Lord GOD unto the shepherds; Woe be to the shepherds of Israel that do feed themselves! should not the shepherds feed the flocks?

hominis propheta de pastoribus Israhel propheta et dices pastoribus haec dicit Dominus Deus uae pastoribus Israhel qui pascebant semet ipsos nonne greges pascuntur a pastoribus 34:3 lac comedebatis et lanis operiebamini et quod crassum 34:3 Ye eat the fat, and ye clothe you with the wool, ye

kill them that are fed: but ye feed not the flock.

erat occidebatis gregem autem meum non pascebatis 34:4 34:4 The diseased have ye not strengthened, neither have ye healed that which was sick, neither have ye bound up that which was broken, neither have ye brought again that which was driven away, neither have ye sought that which was lost; but with force and with cruelty have ye ruled them.

quod infirmum fuit non consolidastis et quod aegrotum non sanastis quod fractum est non alligastis et quod abiectum est non reduxistis quod perierat non quaesistis sed cum austeritate imperabatis eis et cum potentia 34:5 et dispersae sunt 34:5 And they were scattered, because there is no shep-

herd: and they became meat to all the beasts of the field, when they were scattered.

oues meae eo quod non esset pastor et factae sunt in deuorationem omnium bestiarum agri et dispersae sunt 34:6 er- 34:6 My sheep wandered through all the mountains, and

upon every high hill: yea, my flock was scattered upon all the face of the earth, and none did search or seek after

rauerunt greges mei in cunctis montibus et in uniuerso colle them. excelso et super omnem faciem terrae dispersi sunt greges mei et non erat qui requireret non erat inquam qui requireret 34:7 propterea pastores audite uerbum Domini 34:8 34:7 Therefore, ye shepherds, hear the word of the LORD;

34:8 As I live, saith the Lord GOD, surely because my flock became a prey, and my flock became meat to every beast of the field, because there was no shepherd, neither did my shepherds search for my flock, but the shepherds fed themselves, and fed not my flock;

uiuo ego dicit Dominus Deus quia pro eo quod facti sunt greges mei in rapinam et oues meae in deuorationem omnium bestiarum agri eo quod non esset pastor neque enim quaesierunt pastores gregem meum sed pascebant pastores semet ipsos et greges meos non pascebant 34:9 propterea 34:9 Therefore, O ye shepherds, hear the word of the

LORD;

pastores audite uerbum Domini 34:10 haec dicit Dominus 34:10 Thus saith the Lord GOD; Behold, I am against the

shepherds; and I will require my flock at their hand, and cause them to cease from feeding the flock; neither shall the shepherds feed themselves any more; for I will deliver my flock from their mouth, that they may not be meat for them.

Deus ecce ego ipse super pastores requiram gregem meum de manu eorum et cessare eos faciam ut ultra non pascant gregem nec pascant amplius pastores semet ipsos et liberabo gregem meum de ore eorum et non erunt ultra eis in

escam 34:11 quia haec dicit Dominus Deus ecce ego ipse 34:11 For thus saith the Lord GOD; Behold, I, even I, will

both search my sheep, and seek them out.

requiram oues meas et uisitabo eas 34:12 sicut uisitat pas- 34:12 As a shepherd seeketh out his flock in the day that

he is among his sheep that are scattered; so will I seek out my sheep, and will deliver them out of all places where

tor gregem suum in die quando fuerit in medio ouium sua- they have been scattered in the cloudy and dark day.

Biblia Sacra Vulgata

rum dissipatarum sic uisitabo oues meas et liberabo eas de omnibus locis quo dispersae fuerant in die nubis et caliginis

34:13 And I will bring them out from the people, and 34:13 et educam eas de populis et congregabo eas de terris

gather them from the countries, and will bring them to their own land, and feed them upon the mountains of Israel by the rivers, and in all the inhabited places of the country. et inducam eas in terram suam et pascam eas in montibus

34:14 I will feed them in a good pasture, and upon the high Israhel in riuis et in cunctis sedibus terrae 34:14 in pascuis

mountains of Israel shall their fold be: there shall they lie in a good fold, and in a fat pasture shall they feed upon the

mountains of Israel. uberrimis pascam eas et in montibus excelsis Israhel erunt

pascuae eorum ibi requiescent in herbis uirentibus et in pa-

34:15 I will feed my flock, and I will cause them to lie scuis pinguibus pascentur super montes Israhel 34:15 ego

down, saith the Lord GOD.

pascam oues meas et ego eas accubare faciam dicit Domi-

34:16 I will seek that which was lost, and bring again that nus Deus 34:16 quod perierat requiram et quod abiectum which was driven away, and will bind up that which was broken, and will strengthen that which was sick: but I will destroy the fat and the strong; I will feed them with judgment.

erat reducam et quod confractum fuerat alligabo et quod infirmum erat consolidabo et quod pingue et forte custodiam et

34:17 And as for you, O my flock, thus saith the Lord pascam illas in iudicio 34:17 uos autem greges mei haec di-

GOD; Behold, I judge between cattle and cattle, between the rams and the he goats.

cit Dominus Deus ecce ego iudico inter pecus et pecus arie-

34:18 Seemeth it a small thing unto you to have eaten up tum et hircorum 34:18 nonne satis uobis erat pascuam bo-

the good pasture, but ye must tread down with your feet the residue of your pastures? and to have drunk of the

deep waters, but ye must foul the residue with your feet? nam depasci insuper et reliquias pascuarum uestrarum con-

culcastis pedibus uestris et cum purissimam aquam bibere-

34:19 And as for my flock, they eat that which ye have tis reliquam pedibus uestris turbabatis 34:19 et oues meae

trodden with your feet; and they drink that which ye have fouled with your feet.

his quae conculcata pedibus uestris fuerant pascebantur et

34:20 Therefore thus saith the Lord GOD unto them; Be- quae pedes uestri turbauerant haec bibebant 34:20 propte-

hold, I, even I, will judge between the fat cattle and between the lean cattle.

rea haec dicit Dominus Deus ad eos ecce ego ipse iudico

34:21 Because ye have thrust with side and with shoulder, inter pecus pingue et macilentum 34:21 pro eo quod lateand pushed all the diseased with your horns, till ye have scattered them abroad;

ribus et umeris inpingebatis et cornibus uestris uentilabatis

34:22 Therefore will I save my flock, and they shall no omnia infirma pecora donec dispergerentur foras 34:22 sal-

more be a prey; and I will judge between cattle and cattle.

uabo gregem meum et non erit ultra in rapinam et iudicabo

34:23 And I will set up one shepherd over them, and he inter pecus et pecus 34:23 et suscitabo super ea pastorem

shall feed them, even my servant David; he shall feed them, and he shall be their shepherd.

unum qui pascat ea seruum meum Dauid ipse pascet ea et

34:24 And I the LORD will be their God, and my servant ipse erit eis in pastorem 34:24 ego autem Dominus ero eis

David a prince among them; I the LORD have spoken it.

in Deum et seruus meus Dauid princeps in medio eorum ego

34:25 And I will make with them a covenant of peace, and Dominus locutus sum 34:25 et faciam cum eis pactum pawill cause the evil beasts to cease out of the land: and they shall dwell safely in the wilderness, and sleep in the

woods. cis et cessare faciam bestias pessimas de terra et qui habitant 34:26 And I will make them and the places round about in deserto securi dormient in saltibus 34:26 et ponam eos

my hill a blessing; and I will cause the shower to come down in his season; there shall be showers of blessing.

in circuitu collis mei benedictionem et deducam imbrem in

Prophetia Ezechielis

tempore suo pluuiae benedictionis erunt 34:27 et dabit lig- 34:27 And the tree of the field shall yield her fruit, and

the earth shall yield her increase, and they shall be safe in their land, and shall know that I am the LORD, when I have broken the bands of their yoke, and delivered them out of the hand of those that served themselves of them.

num agri fructum suum et terra dabit germen suum et erunt in terra sua absque timore et scient quia ego Dominus cum contriuero catenas iugi eorum et eruero eos de manu imperantium sibi 34:28 et non erunt ultra in rapinam gentibus 34:28 And they shall no more be a prey to the heathen,

neither shall the beast of the land devour them; but they shall dwell safely, and none shall make them afraid.

neque bestiae terrae deuorabunt eos sed habitabunt confidenter absque ullo terrore 34:29 et suscitabo eis germen nomi- 34:29 And I will raise up for them a plant of renown, and

they shall be no more consumed with hunger in the land, neither bear the shame of the heathen any more.

natum et non erunt ultra inminuti fame in terra neque portabunt amplius obprobria gentium 34:30 et scient quia ego 34:30 Thus shall they know that I the LORD their God am

with them, and that they, even the house of Israel, are my people, saith the Lord GOD.

Dominus Deus eorum cum eis et ipsi populus meus domus Israhel ait Dominus Deus 34:31 uos autem greges mei gre- 34:31 And ye my flock, the flock of my pasture, are men,

and I am your God, saith the Lord GOD.

ges pascuae meae homines estis et ego Dominus Deus uester dicit Dominus Deus

35:1 et factus est sermo Domini ad me dicens 35:2 fili 35:1 Moreover the word of the LORD came unto me, say-

ing,

35:2 Son of man, set thy face against mount Seir, and prophesy against it,

hominis pone faciem tuam aduersum montem Seir et prophetabis de eo et dices illi 35:3 haec dicit Dominus Deus 35:3 And say unto it, Thus saith the Lord GOD; Behold, O

mount Seir, I am against thee, and I will stretch out mine hand against thee, and I will make thee most desolate.

ecce ego ad te mons Seir et extendam manum meam super

te et dabo te desolatum atque desertum 35:4 urbes tuas de- 35:4 I will lay thy cities waste, and thou shalt be desolate,

and thou shalt know that I am the LORD.

moliar et tu desertus eris et scies quia ego Dominus 35:5 eo 35:5 Because thou hast had a perpetual hatred, and hast

shed the blood of the children of Israel by the force of the sword in the time of their calamity, in the time that their

quod fueris inimicus sempiternus et concluseris filios Israhel iniquity had an end: in manus gladii in tempore adflictionis eorum in tempore iniquitatis extremae 35:6 propterea uiuo ego dicit Dominus 35:6 Therefore, as I live, saith the Lord GOD, I will pre-

pare thee unto blood, and blood shall pursue thee: sith thou hast not hated blood, even blood shall pursue thee.

Deus quoniam sanguini tradam te et sanguis te persequetur et cum sanguinem oderis sanguis persequetur te 35:7 et 35:7 Thus will I make mount Seir most desolate, and cut

off from it him that passeth out and him that returneth.

dabo montem Seir desolatum et desertum et auferam de eo euntem et redeentem 35:8 et implebo montes eius occiso- 35:8 And I will fill his mountains with his slain men: in

thy hills, and in thy valleys, and in all thy rivers, shall they fall that are slain with the sword.

rum suorum in collibus tuis et in uallibus tuis atque in torrentibus interfecti gladio cadent 35:9 in solitudines sempi- 35:9 I will make thee perpetual desolations, and thy cities

shall not return: and ye shall know that I am the LORD.

ternas tradam te et ciuitates tuae non habitabuntur et scietis quoniam ego Dominus 35:10 eo quod dixeris duae gentes 35:10 Because thou hast said, These two nations and

these two countries shall be mine, and we will possess it; whereas the LORD was there:

et duae terrae meae erunt et hereditate possidebo eas cum Dominus esset ibi 35:11 propterea uiuo ego dicit Dominus 35:11 Therefore, as I live, saith the Lord GOD, I will even

do according to thine anger, and according to thine envy which thou hast used out of thy hatred against them; and I will make myself known among them, when I have judged thee.

Deus quia faciam iuxta iram tuam et secundum zelum tuum

Biblia Sacra Vulgata

quem fecisti odio habens eos et notus efficiar per eos cum

35:12 And thou shalt know that I am the LORD, and that te iudicauero 35:12 et scies quia ego Dominus audiui uni-

I have heard all thy blasphemies which thou hast spoken against the mountains of Israel, saying, They are laid des-

olate, they are given us to consume. uersa obprobria tua quae locutus es de montibus Israhel di-35:13 Thus with your mouth ye have boasted against me, cens deserti nobis dati sunt ad deuorandum 35:13 et insur-

and have multiplied your words against me: I have heard them.

rexistis super me ore uestro et rogastis aduersum me uerba

35:14 Thus saith the Lord GOD; When the whole earth uestra ego audiui 35:14 haec dicit Dominus Deus laetante

rejoiceth, I will make thee desolate.

35:15 As thou didst rejoice at the inheritance of the house uniuersa terra in solitudinem te rediiam 35:15 sicuti gauisus

of Israel, because it was desolate, so will I do unto thee: thou shalt be desolate, O mount Seir, and all Idumea, even

all of it: and they shall know that I am the LORD. es super hereditatem domus Israhel eo quod fuerit dissipata sic faciam tibi dissipatus eris mons Seir et Idumea omnis et scient quia ego Dominus

36:1 Also, thou son of man, prophesy unto the mountains 36:1 tu autem fili hominis propheta super montes Israhel et

of Israel, and say, Ye mountains of Israel, hear the word of the LORD:

dices montes Israhel audite uerbum Domini 36:2 haec dicit

36:2 Thus saith the Lord GOD; Because the enemy hath said against you, Aha, even the ancient high places are ours in possession:

Dominus Deus eo quod dixerit inimicus de uobis euge al-

36:3 Therefore prophesy and say, Thus saith the Lord titudines sempiternae in hereditatem datae sunt nobis 36:3

GOD; Because they have made you desolate, and swallowed you up on every side, that ye might be a possession unto the residue of the heathen, and ye are taken up in the lips of talkers, and are an infamy of the people:

propterea uaticinare et dic haec dicit Dominus Deus pro eo quod desolati estis et conculcati per circuitum et facti in hereditatem reliquis gentibus et ascendistis super labium lin-

36:4 Therefore, ye mountains of Israel, hear the word of guae et obprobrium populi 36:4 propterea montes Israhel

the Lord GOD; Thus saith the Lord GOD to the mountains, and to the hills, to the rivers, and to the valleys, to the desolate wastes, and to the cities that are forsaken, which became a prey and derision to the residue of the heathen that are round about;

audite uerbum Domini Dei haec dicit Dominus Deus montibus et collibus torrentibus uallibusque et desertis parietinis et urbibus derelictis quae depopulatae sunt et subsannatae

36:5 Therefore thus saith the Lord GOD; Surely in the fire a reliquis gentibus per circuitum 36:5 propterea haec dicit

of my jealousy have I spoken against the residue of the heathen, and against all Idumea, which have appointed my land into their possession with the joy of all their heart, with despiteful minds, to cast it out for a prey.

Dominus Deus quoniam in igne zeli mei locutus sum de reliquis gentibus et de Idumea uniuersa qui dederunt terram meam sibi in hereditatem cum gaudio et toto corde ex animo

36:6 Prophesy therefore concerning the land of Israel, and et eiecerunt eam ut uastarent 36:6 idcirco uaticinare super

say unto the mountains, and to the hills, to the rivers, and to the valleys, Thus saith the Lord GOD; Behold, I have spoken in my jealousy and in my fury, because ye have borne the shame of the heathen:

humum Israhel et dices montibus et collibus iugis et uallibus haec dicit Dominus Deus ecce ego in zelo meo et in furore meo locutus sum eo quod confusionem gentium sustinueri-

36:7 Therefore thus saith the Lord GOD; I have lifted up tis 36:7 idcirco haec dicit Dominus Deus ego leuaui mamine hand, Surely the heathen that are about you, they shall bear their shame.

num meam ut gentes quae in circuitu uestro sunt ipsae con-

36:8 But ye, O mountains of Israel, ye shall shoot forth fusionem suam portent 36:8 uos autem montes Israhel ra-

your branches, and yield your fruit to my people of Israel; for they are at hand to come.

mos uestros germinetis et fructum uestrum adferatis populo

Prophetia Ezechielis

meo Israhel prope est enim ut ueniat 36:9 quia ecce ego 36:9 For, behold, I am for you, and I will turn unto you,

and ye shall be tilled and sown:

ad uos et conuertar ad uos et arabimini et accipietis sementem 36:10 et multiplicabo in uobis homines omnemque do- 36:10 And I will multiply men upon you, all the house of

Israel, even all of it: and the cities shall be inhabited, and the wastes shall be builded:

mum Israhel et habitabuntur ciuitates et ruinosa instaurabuntur 36:11 et replebo uos hominibus et iumentis et multipli- 36:11 And I will multiply upon you man and beast; and

they shall increase and bring fruit: and I will settle you after your old estates, and will do better unto you than at

cabuntur et crescent et habitari uos faciam sicut a principio your beginnings: and ye shall know that I am the LORD. bonisque donabo maioribus quam habuistis ab initio et scietis quia ego Dominus 36:12 et adducam super uos homines 36:12 Yea, I will cause men to walk upon you, even my

people Israel; and they shall possess thee, and thou shalt be their inheritance, and thou shalt no more henceforth

populum meum Israhel et hereditate possidebunt te et eris bereave them of men. eis in hereditatem et non addes ultra ut absque eis sis 36:13 36:13 Thus saith the Lord GOD; Because they say unto

you, Thou land devourest up men, and hast bereaved thy nations:

haec dicit Dominus Deus pro eo quod dicunt de uobis deuoratrix hominum es et suffocans gentem tuam 36:14 pro- 36:14 Therefore thou shalt devour men no more, neither

bereave thy nations any more, saith the Lord GOD.

pterea homines non comedes amplius et gentem tuam non necabis ultra ait Dominus Deus 36:15 nec auditam faciam 36:15 Neither will I cause men to hear in thee the shame

of the heathen any more, neither shalt thou bear the reproach of the people any more, neither shalt thou cause

in te amplius confusionem gentium et obprobrium populo- thy nations to fall any more, saith the Lord GOD. rum nequaquam portabis et gentem tuam non amittes amplius ait Dominus Deus 36:16 et factum est uerbum Domini 36:16 Moreover the word of the LORD came unto me,

saying,

ad me dicens 36:17 fili hominis domus Israhel habitauerunt 36:17 Son of man, when the house of Israel dwelt in their

own land, they defiled it by their own way and by their doings: their way was before me as the uncleanness of a

in humo sua et polluerunt eam in uiis suis et in studiis suis removed woman. iuxta inmunditiam menstruatae facta est uia eorum coram me 36:18 et effudi indignationem meam super eos pro san- 36:18 Wherefore I poured my fury upon them for the

blood that they had shed upon the land, and for their idols wherewith they had polluted it:

guine quem fuderunt super terram et in idolis suis polluerunt eam 36:19 et dispersi eos in gentes et uentilati sunt in terris 36:19 And I scattered them among the heathen, and they

were dispersed through the countries: according to their way and according to their doings I judged them.

iuxta uias eorum et adinuentiones iudicaui eos 36:20 et in-

36:20 And when they entered unto the heathen, whither they went, they profaned my holy name, when they said to them, These are the people of the LORD, and are gone forth out of his land.

gressi sunt ad gentes ad quas introierunt et polluerunt nomen sanctum meum cum diceretur de eis populus Domini iste est et de terra eius egressi sunt 36:21 et peperci nomini meo 36:21 But I had pity for mine holy name, which the house

of Israel had profaned among the heathen, whither they went.

sancto quod polluerat domus Israhel in gentibus ad quas ingressi sunt 36:22 idcirco dices domui Israhel haec dicit Do- 36:22 Therefore say unto the house of Israel, thus saith the

Lord GOD; I do not this for your sakes, O house of Israel, but for mine holy name’s sake, which ye have profaned

minus Deus non propter uos ego faciam domus Israhel sed among the heathen, whither ye went. propter nomen sanctum meum quod polluistis in gentibus ad quas intrastis 36:23 et sanctificabo nomen meum mag- 36:23 And I will sanctify my great name, which was pro-

faned among the heathen, which ye have profaned in the midst of them; and the heathen shall know that I am the LORD, saith the Lord GOD, when I shall be sanctified in you before their eyes.

num quod pollutum est inter gentes quod polluistis in medio

Biblia Sacra Vulgata

earum ut sciant gentes quia ego Dominus ait Dominus ex-

36:24 For I will take you from among the heathen, and ercituum cum sanctificatus fuero in uobis coram eis 36:24

gather you out of all countries, and will bring you into your own land.

tollam quippe uos de gentibus et congregabo de uniuersis

36:25 Then will I sprinkle clean water upon you, and ye terris et adducam uos in terram uestram 36:25 et effundam

shall be clean: from all your filthiness, and from all your idols, will I cleanse you.

super uos aquam mundam et mundabimini ab omnibus inquinamentis uestris et ab uniuersis idolis uestris mundabo

36:26 A new heart also will I give you, and a new spirit uos 36:26 et dabo uobis cor nouum et spiritum nouum po-

will I put within you: and I will take away the stony heart out of your flesh, and I will give you an heart of flesh.

nam in medio uestri et auferam cor lapideum de carne uestra

36:27 And I will put my spirit within you, and cause you et dabo uobis cor carneum 36:27 et spiritum meum ponam to walk in my statutes, and ye shall keep my judgments, and do them.

in medio uestri et faciam ut in praeceptis meis ambuletis et

36:28 And ye shall dwell in the land that I gave to your iudicia mea custodiatis et operemini 36:28 et habitabitis in

fathers; and ye shall be my people, and I will be your God.

terra quam dedi patribus uestris et eritis mihi in populum

36:29 I will also save you from all your uncleannesses: et ego ero uobis in Deum 36:29 et saluabo uos ex uniuerand I will call for the corn, and will increase it, and lay no famine upon you.

sis inquinamentis uestris et uocabo frumentum et multipli-

36:30 And I will multiply the fruit of the tree, and the cabo illud et non inponam in uobis famem 36:30 et multi-

increase of the field, that ye shall receive no more reproach of famine among the heathen.

plicabo fructum ligni et genimina agri ut non portetis ultra

36:31 Then shall ye remember your own evil ways, and obprobrium famis in gentibus 36:31 et recordabimini uia-

your doings that were not good, and shall lothe yourselves in your own sight for your iniquities and for your abomi-

nations. rum uestrarum pessimarum studiorumque non bonorum et 36:32 Not for your sakes do I this, saith the Lord GOD, be displicebunt uobis iniquitates uestrae et scelera uestra 36:32

it known unto you: be ashamed and confounded for your own ways, O house of Israel.

non propter uos ego faciam ait Dominus Deus notum sit uobis confundimini et erubescite super uiis uestris domus Isra-

36:33 Thus saith the Lord GOD; In the day that I shall hel 36:33 haec dicit Dominus Deus in die qua mundauero

have cleansed you from all your iniquities I will also cause you to dwell in the cities, and the wastes shall be builded.

uos ex omnibus iniquitatibus uestris et habitari fecero urbes

36:34 And the desolate land shall be tilled, whereas it lay et instaurauero ruinosa 36:34 et terra deserta fuerit exculta

desolate in the sight of all that passed by. 36:35 And they shall say, This land that was desolate is be- quae quondam erat desolata in oculis omnis uiatoris 36:35 come like the garden of Eden; and the waste and desolate and ruined cities are become fenced, and are inhabited.

dicent terra illa inculta facta est ut hortus uoluptatis et ciuitates desertae et destitutae atque suffossae munitae sederunt

36:36 Then the heathen that are left round about you shall 36:36 et scient gentes quaecumque derelictae fuerint in cir-

know that I the LORD build the ruined places, and plant that that was desolate: I the LORD have spoken it, and I

will do it. cuitu uestro quia ego Dominus aedificaui dissipata plantaui-36:37 Thus saith the Lord GOD; I will yet for this be en- que inculta ego Dominus locutus sum et fecerim 36:37 haec

quired of by the house of Israel, to do it for them; I will increase them with men like a flock.

dicit Dominus Deus adhuc in hoc inuenient me domus Israhel ut faciam eis multiplicabo eos sicut gregem hominum

36:38 As the holy flock, as the flock of Jerusalem in her 36:38 ut gregem sanctum ut gregem Hierusalem in sollem-

solemn feasts; so shall the waste cities be filled with flocks of men: and they shall know that I am the LORD.

Prophetia Ezechielis

nitatibus eius sic erunt ciuitates desertae plenaeque gregibus hominum et scient quia ego Dominus

37:1 facta est super me manus Domini et eduxit me in spi- 37:1 The hand of the LORD was upon me, and carried

me out in the spirit of the LORD, and set me down in the midst of the valley which was full of bones,

ritu Domini et dimisit me in medio campi qui erat plenus ossibus 37:2 et circumduxit me per ea in gyro erant autem 37:2 And caused me to pass by them round about: and,

behold, there were very many in the open valley; and, lo, they were very dry.

multa ualde super faciem campi siccaque uehementer 37:3

37:3 And he said unto me, Son of man, can these bones live? And I answered, O Lord GOD, thou knowest.

et dixit ad me fili hominis putasne uiuent ossa ista et dixi Domine Deus tu nosti 37:4 et dixit ad me uaticinare de ossibus 37:4 Again he said unto me, Prophesy upon these bones,

and say unto them, O ye dry bones, hear the word of the LORD.

istis et dices eis ossa arida audite uerbum Domini 37:5 haec

37:5 Thus saith the Lord GOD unto these bones; Behold, I will cause breath to enter into you, and ye shall live:

dicit Dominus Deus ossibus his ecce ego intromittam in uos spiritum et uiuetis 37:6 et dabo super uos neruos et succ- 37:6 And I will lay sinews upon you, and will bring up

flesh upon you, and cover you with skin, and put breath in you, and ye shall live; and ye shall know that I am the

rescere faciam super uos carnes et superextendam in uobis LORD. cutem et dabo uobis spiritum et uiuetis et scietis quia ego Dominus 37:7 et prophetaui sicut praeceperat mihi factus 37:7 So I prophesied as I was commanded: and as I proph-

esied, there was a noise, and behold a shaking, and the bones came together, bone to his bone.

est autem sonitus prophetante me et ecce commotio et accesserunt ossa ad ossa unumquodque ad iuncturam suam 37:8 37:8 And when I beheld, lo, the sinews and the flesh came

up upon them, and the skin covered them above: but there was no breath in them.

et uidi et ecce super ea nerui et carnes ascenderunt et extenta est in eis cutis desuper et spiritum non habebant 37:9 37:9 Then said he unto me, Prophesy unto the wind,

prophesy, son of man, and say to the wind, Thus saith the Lord GOD; Come from the four winds, O breath, and

et dixit ad me uaticinare ad spiritum uaticinare fili hominis et breathe upon these slain, that they may live. dices ad spiritum haec dicit Dominus Deus a quattuor uentis ueni spiritus et insufla super interfectos istos et reuiuescant 37:10 et prophetaui sicut praeceperat mihi et ingressus est 37:10 So I prophesied as he commanded me, and the

breath came into them, and they lived, and stood up upon their feet, an exceeding great army.

in ea spiritus et uixerunt steteruntque super pedes suos exercitus grandis nimis ualde 37:11 et dixit ad me fili homi- 37:11 Then he said unto me, Son of man, these bones are

the whole house of Israel: behold, they say, Our bones are dried, and our hope is lost: we are cut off for our parts.

nis ossa haec uniuersa domus Israhel est ipsi dicunt aruerunt ossa nostra et periit spes nostra et abscisi sumus 37:12 pro- 37:12 Therefore prophesy and say unto them, Thus saith

the Lord GOD; Behold, O my people, I will open your graves, and cause you to come up out of your graves, and

pterea uaticinare et dices ad eos haec dicit Dominus Deus bring you into the land of Israel. ecce ego aperiam tumulos uestros et educam uos de sepulchris uestris populus meus et inducam uos in terram Israhel 37:13 et scietis quia ego Dominus cum aperuero sepulchra 37:13 And ye shall know that I am the LORD, when I have

opened your graves, O my people, and brought you up out of your graves,

uestra et eduxero uos de tumulis uestris populus meus 37:14

37:14 And shall put my spirit in you, and ye shall live, and I shall place you in your own land: then shall ye know that I the LORD have spoken it, and performed it, saith the LORD.

et dedero spiritum meum in uobis et uixeritis et requiescere uos faciam super humum uestram et scietis quia ego Domi-

Biblia Sacra Vulgata

37:15 The word of the LORD came again unto me, saying, nus locutus sum et feci ait Dominus Deus 37:15 et factus est

37:16 Moreover, thou son of man, take thee one stick, and sermo Domini ad me dicens 37:16 et tu fili hominis sume write upon it, For Judah, and for the children of Israel his companions: then take another stick, and write upon it, For Joseph, the stick of Ephraim and for all the house of Israel his companions:

tibi lignum unum et scribe super illud Iudae et filiorum Israhel sociis eius et tolle lignum alterum et scribe super eum Ioseph lignum Ephraim et cunctae domui Israhel sociorum-

37:17 And join them one to another into one stick; and que eius 37:17 et adiunge illa unum ad alterum tibi in li-

they shall become one in thine hand. 37:18 And when the children of thy people shall speak gnum unum et erunt in unionem in manu tua 37:18 cum unto thee, saying, Wilt thou not shew us what thou meanest by these?

autem dixerint ad te filii populi tui loquentes nonne indicas

37:19 Say unto them, Thus saith the Lord GOD; Behold, nobis quid in his tibi uelis 37:19 loqueris ad eos haec di-

I will take the stick of Joseph, which is in the hand of Ephraim, and the tribes of Israel his fellows, and will put them with him, even with the stick of Judah, and make them one stick, and they shall be one in mine hand.

cit Dominus Deus ecce ego adsumam lignum Ioseph quod est in manu Ephraim et tribus Israhel quae iunctae sunt ei et dabo eas pariter cum ligno Iuda et faciam eas in lignum

37:20 And the sticks whereon thou writest shall be in thine unum et erunt unum in manu eius 37:20 erunt autem ligna hand before their eyes.

37:21 And say unto them, Thus saith the Lord GOD; Be- super quae scripseris in manu tua in oculis eorum 37:21 et

hold, I will take the children of Israel from among the heathen, whither they be gone, and will gather them on every

side, and bring them into their own land: dices ad eos haec dicit Dominus Deus ecce ego adsumam

filios Israhel de medio nationum ad quas abierunt et congre-

37:22 And I will make them one nation in the land upon gabo eos undique et adducam eos ad humum suam 37:22 et

the mountains of Israel; and one king shall be king to them all: and they shall be no more two nations, neither shall

they be divided into two kingdoms any more at all. faciam eos gentem unam in terra in montibus Israhel et rex

unus erit omnibus imperans et non erunt ultra duae gentes

37:23 Neither shall they defile themselves any more with nec diuidentur amplius in duo regna 37:23 neque polluen-

their idols, nor with their detestable things, nor with any of their transgressions: but I will save them out of all their dwellingplaces, wherein they have sinned, and will cleanse them: so shall they be my people, and I will be their God.

tur ultra in idolis suis et abominationibus suis et in cunctis iniquitatibus suis et saluos eos faciam de uniuersis sedibus suis in quibus peccauerunt et mundabo eos et erunt mihi po-

37:24 And David my servant shall be king over them; and pulus et ego ero eis Deus 37:24 et seruus meus Dauid rex they all shall have one shepherd: they shall also walk in my judgments, and observe my statutes, and do them.

super eos et pastor unus erit omnium eorum in iudiciis meis

37:25 And they shall dwell in the land that I have given ambulabunt et mandata mea custodient et facient ea 37:25

unto Jacob my servant, wherein your fathers have dwelt; and they shall dwell therein, even they, and their children, and their children’s children for ever: and my servant David shall be their prince for ever.

et habitabunt super terram quam dedi seruo meo Iacob in qua habitauerunt patres uestri et habitabunt super eam ipsi et filii eorum et filii filiorum eorum usque in sempiternum et

37:26 Moreover I will make a covenant of peace with Dauid seruus meus princeps eorum in perpetuum 37:26 et

them; it shall be an everlasting covenant with them: and I will place them, and multiply them, and will set my sanc-

tuary in the midst of them for evermore. percutiam illis foedus pacis pactum sempiternum erit eis et

fundabo eos et multiplicabo et dabo sanctificationem meam

37:27 My tabernacle also shall be with them: yea, I will in medio eorum in perpetuum 37:27 et erit tabernaculum be their God, and they shall be my people.

Prophetia Ezechielis

meum in eis et ero eis Deus et ipsi erunt mihi populus 37:28 37:28 And the heathen shall know that I the LORD do

sanctify Israel, when my sanctuary shall be in the midst of them for evermore.

et scient gentes quia ego Dominus sanctificator Israhel cum fuerit sanctificatio mea in medio eorum in perpetuum

38:1 et factus est sermo Domini ad me dicens 38:2 fili 38:1 And the word of the LORD came unto me, saying,

38:2 Son of man, set thy face against Gog, the land of Magog, the chief prince of Meshech and Tubal, and prophesy against him,

hominis pone faciem tuam contra Gog terram Magog principem capitis Mosoch et Thubal et uaticinare de eo 38:3 38:3 And say, Thus saith the Lord GOD; Behold, I am

against thee, O Gog, the chief prince of Meshech and Tubal:

et dices ad eum haec dicit Dominus Deus ecce ego ad te

Gog principem capitis Mosoch et Thubal 38:4 et circuma- 38:4 And I will turn thee back, and put hooks into thy jaws, and I will bring thee forth, and all thine army, horses and horsemen, all of them clothed with all sorts of armour, even a great company with bucklers and shields, all of them handling swords:

gam te et ponam frenum in maxillis tuis et educam te et omnem exercitum tuum equos et equites uestitos loricis uniuersos multitudinem magnam hastam et clypeum arripientium

et gladium 38:5 Persae Aethiopes et Lybies cum eis om- 38:5 Persia, Ethiopia, and Libya with them; all of them

with shield and helmet:

nes scutati et galeati 38:6 Gomer et uniuersa agmina eius 38:6 Gomer, and all his bands; the house of Togarmah of

the north quarters, and all his bands: and many people with thee.

domus Thogorma latera aquilonis et totum robur eius po-

pulique multi tecum 38:7 praepara et instrue te et omnem 38:7 Be thou prepared, and prepare for thyself, thou, and all thy company that are assembled unto thee, and be thou a guard unto them.

multitudinem tuam quae coaceruata est ad te et esto eis in praeceptum 38:8 post dies multos uisitaberis in nouissimo 38:8 After many days thou shalt be visited: in the lat-

ter years thou shalt come into the land that is brought back from the sword, and is gathered out of many people, against the mountains of Israel, which have been always waste: but it is brought forth out of the nations, and they shall dwell safely all of them.

annorum uenies ad terram quae reuersa est a gladio congregata est de populis multis ad montes Israhel qui fuerunt deserti iugiter haec de populis educta est et habitauerunt in ea confidenter uniuersi 38:9 ascendens autem quasi tempestas 38:9 Thou shalt ascend and come like a storm, thou shalt

be like a cloud to cover the land, thou, and all thy bands, and many people with thee.

uenies et quasi nubes ut operias terram tu et omnia agmina tua et populi multi tecum 38:10 haec dicit Dominus Deus in 38:10 Thus saith the Lord GOD; It shall also come to pass,

that at the same time shall things come into thy mind, and thou shalt think an evil thought:

die illa ascendent sermones super cor tuum et cogitabis cogitationem pessimam 38:11 et dices ascendam ad terram abs- 38:11 And thou shalt say, I will go up to the land of un-

walled villages; I will go to them that are at rest, that dwell safely, all of them dwelling without walls, and having nei-

que muro ueniam ad quiescentes habitantesque secure om- ther bars nor gates, nes habitant sine muro uectes et portae non sunt eis 38:12 38:12 To take a spoil, and to take a prey; to turn thine hand

upon the desolate places that are now inhabited, and upon the people that are gathered out of the nations, which have

ut diripias spolia et inuadas praedam ut inferas manum tuam gotten cattle and goods, that dwell in the midst of the land. super eos qui deserti fuerant et postea restituti et super populum qui est congregatus ex gentibus qui possidere coepit et esse habitator umbilici terrae 38:13 Seba et Dedan et ne- 38:13 Sheba, and Dedan, and the merchants of Tarshish,

with all the young lions thereof, shall say unto thee, Art thou come to take a spoil? hast thou gathered thy company to take a prey? to carry away silver and gold, to take away cattle and goods, to take a great spoil?

gotiatores Tharsis et omnes leones eius dicent tibi numquid ad sumenda spolia tu uenis ecce ad diripiendam praedam

Biblia Sacra Vulgata

congregasti multitudinem tuam ut tollas argentum et aurum auferas supellectilem atque substantiam et diripias manubias

38:14 Therefore, son of man, prophesy and say unto Gog, infinitas 38:14 propterea uaticinare fili hominis et dices ad

Thus saith the Lord GOD; In that day when my people of Israel dwelleth safely, shalt thou not know it?

Gog haec dicit Dominus Deus numquid non in die illo cum

38:15 And thou shalt come from thy place out of the north habitauerit populus meus Israhel confidenter scies 38:15 et

parts, thou, and many people with thee, all of them riding upon horses, a great company, and a mighty army:

uenies de loco tuo a lateribus aquilonis tu et populi multi tecum ascensores equorum uniuersi coetus magnus et exer-

38:16 And thou shalt come up against my people of Israel, citus uehemens 38:16 et ascendes super populum meum Isas a cloud to cover the land; it shall be in the latter days, and I will bring thee against my land, that the heathen may know me, when I shall be sanctified in thee, O Gog, before their eyes.

rahel quasi nubes ut operias terram in nouissimis diebus eris et adducam te super terram meam ut sciant gentes me cum

38:17 Thus saith the Lord GOD; Art thou he of whom I sanctificatus fuero in te in oculis eorum o Gog 38:17 haec

have spoken in old time by my servants the prophets of Israel, which prophesied in those days many years that I

would bring thee against them? dicit Dominus Deus tu ergo ille es de quo locutus sum in diebus antiquis in manu seruorum meorum prophetarum Israhel qui prophetauerunt in diebus illorum temporum ut adduce-

38:18 And it shall come to pass at the same time when rem te super eos 38:18 et erit in die illa in die aduentus Gog

Gog shall come against the land of Israel, saith the Lord GOD, that my fury shall come up in my face.

super terram Israhel ait Dominus Deus ascendet indignatio

38:19 For in my jealousy and in the fire of my wrath have mea in furore meo 38:19 et in zelo meo in igne irae meae lo-

I spoken, Surely in that day there shall be a great shaking in the land of Israel;

cutus sum quia in die illa erit commotio magna super terram

38:20 So that the fishes of the sea, and the fowls of the Israhel 38:20 et commouebuntur a facie mea pisces maris

heaven, and the beasts of the field, and all creeping things that creep upon the earth, and all the men that are upon the face of the earth, shall shake at my presence, and the mountains shall be thrown down, and the steep places shall fall, and every wall shall fall to the ground.

et uolucres caeli et bestiae agri et omne reptile quod mouetur super humum cunctique homines qui sunt super faciem terrae et subuertentur montes et cadent sepes et omnis murus

38:21 And I will call for a sword against him throughout in terra corruet 38:21 et conuocabo aduersum eum in cunc-

all my mountains, saith the Lord GOD: every man’s sword shall be against his brother.

tis montibus meis gladium ait Dominus Deus gladius uniu-

38:22 And I will plead against him with pestilence and scuiusque in fratrem suum dirigetur 38:22 et iudicabo eum

with blood; and I will rain upon him, and upon his bands, and upon the many people that are with him, an overflow-

ing rain, and great hailstones, fire, and brimstone. peste et sanguine et imbre uehementi et lapidibus inmensis

ignem et sulphur pluam super eum et super exercitum eius et

38:23 Thus will I magnify myself, and sanctify myself; super populos multos qui sunt cum eo 38:23 et magnificaand I will be known in the eyes of many nations, and they shall know that I am the LORD.

bor et sanctificabor et notus ero in oculis gentium multarum et scient quia ego Dominus

39:1 Therefore, thou son of man, prophesy against Gog, 39:1 tu autem fili hominis uaticinare aduersum Gog et di-

and say, Thus saith the Lord GOD; Behold, I am against thee, O Gog, the chief prince of Meshech and Tubal:

ces haec dicit Dominus Deus ecce ego super te Gog princi-

39:2 And I will turn thee back, and leave but the sixth part pem capitis Mosoch et Thubal 39:2 et circumagam te et seof thee, and will cause thee to come up from the north parts, and will bring thee upon the mountains of Israel:

Prophetia Ezechielis

ducam te et ascendere faciam de lateribus aquilonis et adducam te super montes Israhel 39:3 et percutiam arcum tuum 39:3 And I will smite thy bow out of thy left hand, and

will cause thine arrows to fall out of thy right hand.

in manu sinistra tua et sagittas tuas de manu dextera tua deiciam 39:4 super montes Israhel cades tu et omnia agmina 39:4 Thou shalt fall upon the mountains of Israel, thou,

and all thy bands, and the people that is with thee: I will give thee unto the ravenous birds of every sort, and to the

tua et populi qui sunt tecum feris auibus omnique uolatili beasts of the field to be devoured. et bestiis terrae dedi te deuorandum 39:5 super faciem agri 39:5 Thou shalt fall upon the open field: for I have spoken

it, saith the Lord GOD.

cades quia ego locutus sum ait Dominus Deus 39:6 et emit- 39:6 And I will send a fire on Magog, and among them

that dwell carelessly in the isles: and they shall know that I am the LORD.

tam ignem in Magog et in his qui habitant in insulis confidenter et scient quia ego Dominus 39:7 et nomen sanctum 39:7 So will I make my holy name known in the midst of

my people Israel; and I will not let them pollute my holy name any more: and the heathen shall know that I am the

meum notum faciam in medio populi mei Israhel et non pol- LORD, the Holy One in Israel. luam nomen sanctum meum amplius et scient gentes quia ego Dominus Sanctus Israhel 39:8 ecce uenit et factum est 39:8 Behold, it is come, and it is done, saith the Lord

GOD; this is the day whereof I have spoken.

ait Dominus Deus haec est dies de qua locutus sum 39:9 et 39:9 And they that dwell in the cities of Israel shall go

forth, and shall set on fire and burn the weapons, both the shields and the bucklers, the bows and the arrows, and the handstaves, and the spears, and they shall burn them with fire seven years:

egredientur habitatores de ciuitatibus Israhel et succendent et conburent arma clypeum et hastas arcum et sagittas et baculos manus et contos et succendent ea igne septem annis 39:10 et non portabunt ligna de regionibus neque succident 39:10 So that they shall take no wood out of the field, nei-

ther cut down any out of the forests; for they shall burn the weapons with fire: and they shall spoil those that spoiled

de saltibus quoniam arma succendent igne et depraedabun- them, and rob those that robbed them, saith the Lord GOD. tur eos quibus praedae fuerant et diripient uastatores suos ait Dominus Deus 39:11 et erit in die illa dabo Gog locum 39:11 And it shall come to pass in that day, that I will

give unto Gog a place there of graves in Israel, the valley of the passengers on the east of the sea: and it shall stop the noses of the passengers: and there shall they bury Gog and all his multitude: and they shall call it The valley of Hamongog.

nominatum sepulchrum in Israhel uallem Viatorum ad orientem maris quae obstupescere facit praetereuntes et sepelient ibi Gog et omnem multitudinem eius et uocabitur uallis Multitudinis Gog 39:12 et sepelient eos domus Israhel ut mun- 39:12 And seven months shall the house of Israel be bury-

ing of them, that they may cleanse the land.

dent terram septem mensibus 39:13 sepeliet autem omnis 39:13 Yea, all the people of the land shall bury them; and it shall be to them a renown the day that I shall be glorified, saith the Lord GOD.

populus terrae et erit eis nominata dies in qua glorificatus sum ait Dominus Deus 39:14 et uiros iugiter constituent lu- 39:14 And they shall sever out men of continual employ-

ment, passing through the land to bury with the passengers those that remain upon the face of the earth, to cleanse it:

strantes terram qui sepeliant et requirant eos qui remanse- after the end of seven months shall they search. rant super faciem terrae ut emundent eam post menses autem septem quaerere incipient 39:15 et circumibunt perag- 39:15 And the passengers that pass through the land, when

any seeth a man’s bone, then shall he set up a sign by it, till the buriers have buried it in the valley of Hamongog.

rantes terram cumque uiderint os hominis statuent iuxta illud titulum donec sepeliant illud pollinctores in ualle Multitudinis Gog 39:16 nomen autem ciuitatis Amona et mun- 39:16 And also the name of the city shall be Hamonah.

Thus shall they cleanse the land.

Biblia Sacra Vulgata

39:17 And, thou son of man, thus saith the Lord GOD; dabunt terram 39:17 tu ergo fili hominis haec dicit Domi-

Speak unto every feathered fowl, and to every beast of the field, Assemble yourselves, and come; gather yourselves on every side to my sacrifice that I do sacrifice for you, even a great sacrifice upon the mountains of Israel, that ye may eat flesh, and drink blood.

nus Deus dic omni uolucri et uniuersis auibus cunctisque bestiis agri conuenite properate concurrite undique ad uictimam meam quam ego immolo uobis uictimam grandem super montes Israhel ut comedatis carnes et bibatis sanguinem

39:18 Ye shall eat the flesh of the mighty, and drink the 39:18 carnes fortium comedetis et sanguinem principum ter-

blood of the princes of the earth, of rams, of lambs, and of goats, of bullocks, all of them fatlings of Bashan.

rae bibetis arietum agnorum et hircorum taurorumque alti-

39:19 And ye shall eat fat till ye be full, and drink blood lium et pinguium omnium 39:19 et comedetis adipem in satill ye be drunken, of my sacrifice which I have sacrificed for you.

turitate et bibetis sanguinem in ebrietate de uictima quam

39:20 Thus ye shall be filled at my table with horses and ego immolabo uobis 39:20 et saturabimini super mensam

chariots, with mighty men, and with all men of war, saith the Lord GOD.

meam de equo et de equite forti et de uniuersis uiris bella-

39:21 And I will set my glory among the heathen, and all toribus ait Dominus Deus 39:21 et ponam gloriam meam in

the heathen shall see my judgment that I have executed, and my hand that I have laid upon them.

gentibus et uidebunt omnes gentes iudicium meum quod fe-

39:22 So the house of Israel shall know that I am the cerim et manum meam quam posuerim super eos 39:22 et

LORD their God from that day and forward.

scient domus Israhel quia ego Dominus Deus eorum a die

39:23 And the heathen shall know that the house of Israel illa et deinceps 39:23 et scient gentes quoniam in iniqui-

went into captivity for their iniquity: because they trespassed against me, therefore hid I my face from them, and gave them into the hand of their enemies: so fell they all by the sword.

tate sua capta sit domus Israhel eo quod reliquerint me et absconderim faciem meam ab eis et tradiderim eos in manu

39:24 According to their uncleanness and according to hostium et ceciderint in gladio uniuersi 39:24 iuxta inmun-

their transgressions have I done unto them, and hid my face from them.

ditiam eorum et scelus feci eis et abscondi faciem meam ab

39:25 Therefore thus saith the Lord GOD; Now will I illis 39:25 propterea haec dicit Dominus Deus nunc redu-

bring again the captivity of Jacob, and have mercy upon the whole house of Israel, and will be jealous for my holy

name; cam captiuitatem Iacob et miserebor omnis domus Israhel et

39:26 After that they have borne their shame, and all their adsumam zelum pro nomine sancto meo 39:26 et portabunt trespasses whereby they have trespassed against me, when they dwelt safely in their land, and none made them afraid.

confusionem suam et omnem praeuaricationem quam praeuaricati sunt in me cum habitauerint in terra sua confidenter

39:27 When I have brought them again from the people, neminem formidantes 39:27 et reduxero eos de populis et

and gathered them out of their enemies’ lands, and am sanctified in them in the sight of many nations;

congregauero de terris inimicorum suorum et sanctificatus

39:28 Then shall they know that I am the LORD their God, fuero in eis in oculis gentium plurimarum 39:28 et scient

which caused them to be led into captivity among the heathen: but I have gathered them unto their own land, and

have left none of them any more there. quia ego Dominus Deus eorum eo quod transtulerim eos in

nationes et congregauero eos super terram suam et non de-

39:29 Neither will I hide my face any more from them: for reliquerim quemquam ex eis ibi 39:29 et non abscondam I have poured out my spirit upon the house of Israel, saith the Lord GOD.

ultra faciem meam ab eis eo quod effuderim spiritum meum super omnem domum Israhel ait Dominus Deus

Prophetia Ezechielis

40:1 in uicesimo et quinto anno transmigrationis nostrae in 40:1 In the five and twentieth year of our captivity, in the

beginning of the year, in the tenth day of the month, in the fourteenth year after that the city was smitten, in the selfsame day the hand of the LORD was upon me, and brought me thither.

exordio anni decima mensis quartodecimo anno postquam percussa est ciuitas in ipsa hac die facta est super me manus Domini et adduxit me illuc 40:2 in uisionibus Dei ad- 40:2 In the visions of God brought he me into the land of

Israel, and set me upon a very high mountain, by which was as the frame of a city on the south.

duxit me in terram Israhel et dimisit me super montem excelsum nimis super quem erat quasi aedificium ciuitatis uergentis ad austrum 40:3 et introduxit me illuc et ecce uir cuius 40:3 And he brought me thither, and, behold, there was a

man, whose appearance was like the appearance of brass, with a line of flax in his hand, and a measuring reed; and

erat species quasi species aeris et funiculus lineus in manu he stood in the gate. eius et calamus mensurae in manu eius stabat autem in porta 40:4 et locutus est ad me idem uir fili hominis uide oculis 40:4 And the man said unto me, Son of man, behold with

thine eyes, and hear with thine ears, and set thine heart upon all that I shall shew thee; for to the intent that I might shew them unto thee art thou brought hither: declare all that thou seest to the house of Israel.

tuis et auribus tuis audi et pone cor tuum in omnia quae ego ostendam tibi quia ut ostendantur tibi adductus es huc adnuntia omnia quae tu uides domui Israhel 40:5 et ecce mu- 40:5 And behold a wall on the outside of the house round

about, and in the man’s hand a measuring reed of six cubits long by the cubit and an hand breadth: so he measured the

rus forinsecus in circuitu domus undique et in manu uiri ca- breadth of the building, one reed; and the height, one reed. lamus mensurae sex cubitorum et palmo et mensus est latitudinem aedificii calamo uno altitudinem quoque calamo uno 40:6 et uenit ad portam quae respiciebat uiam orienta- 40:6 Then came he unto the gate which looketh toward

the east, and went up the stairs thereof, and measured the threshold of the gate, which was one reed broad; and the

lem et ascendit per gradus eius et mensus est limen portae other threshold of the gate, which was one reed broad. calamo uno latitudinem id est limen unum calamo uno in latitudine 40:7 et thalamum uno calamo in longum et uno 40:7 And every little chamber was one reed long, and one

reed broad; and between the little chambers were five cubits; and the threshold of the gate by the porch of the gate

calamo in latum et inter thalamos quinque cubitos 40:8 et within was one reed.

40:8 He measured also the porch of the gate within, one reed.

limen portae iuxta uestibulum portae intrinsecus calamo uno 40:9 et mensus est uestibulum portae octo cubitorum et fron- 40:9 Then measured he the porch of the gate, eight cubits;

and the posts thereof, two cubits; and the porch of the gate was inward.

tem eius duobus cubitis uestibulum autem portae erat intrinsecus 40:10 porro thalami portae ad uiam orientalem tres 40:10 And the little chambers of the gate eastward were

three on this side, and three on that side; they three were of one measure: and the posts had one measure on this

hinc et tres inde mensura una trium et mensura una fron- side and on that side. tium ex utraque parte 40:11 et mensus est latitudinem li- 40:11 And he measured the breadth of the entry of the

gate, ten cubits; and the length of the gate, thirteen cubits.

minis portae decem cubitorum et longitudinem portae tredecim cubitorum 40:12 et marginem ante thalamos cubiti 40:12 The space also before the little chambers was one

cubit on this side, and the space was one cubit on that side: and the little chambers were six cubits on this side, and six

unius et cubitus unus finis utrimque thalami autem sex cubi- cubits on that side. torum erant hinc et inde 40:13 et mensus est portam a tecto 40:13 He measured then the gate from the roof of one little

chamber to the roof of another: the breadth was five and twenty cubits, door against door.

thalami usque ad tectum eius latitudinem uiginti et quinque cubitorum ostium contra ostium 40:14 et fecit frontes per 40:14 He made also posts of threescore cubits, even unto

the post of the court round about the gate.

Biblia Sacra Vulgata

sexaginta cubitos et ad frontem atrium portae undique per

40:15 And from the face of the gate of the entrance unto circuitum 40:15 et ante faciem portae quae pertingebat us-

the face of the porch of the inner gate were fifty cubits.

que ad faciem uestibuli portae interioris quinquaginta cubi-

40:16 And there were narrow windows to the little cham- tos 40:16 et fenestras obliquas in thalamis et in frontibus

bers, and to their posts within the gate round about, and likewise to the arches: and windows were round about in-

ward: and upon each post were palm trees. eorum quae erant intra portam undique per circuitum simili-

ter autem erant et in uestibulis fenestrae per gyrum intrinse-

40:17 Then brought he me into the outward court, and, lo, cus et ante frontes pictura palmarum 40:17 et eduxit me ad

there were chambers, and a pavement made for the court round about: thirty chambers were upon the pavement.

atrium exterius et ecce gazofilacia et pauimentum stratum lapide in atrio per circuitum triginta gazofilacia in circuitu pa-

40:18 And the pavement by the side of the gates over uimenti 40:18 et pauimentum in fronte portarum secundum

against the length of the gates was the lower pavement.

40:19 Then he measured the breadth from the forefront of longitudinem portarum erat inferius 40:19 et mensus est la-

the lower gate unto the forefront of the inner court without, an hundred cubits eastward and northward.

titudinem a facie portae inferioris usque ad frontem atrii interioris extrinsecus centum cubitos ad orientem et ad aquilo-

40:20 And the gate of the outward court that looked toward nem 40:20 portam quoque quae respiciebat uiam aquilonis the north, he measured the length thereof, and the breadth thereof.

atrii exterioris mensus est tam in longitudine quam in latitu-

40:21 And the little chambers thereof were three on this dine 40:21 et thalamos eius tres hinc et tres inde et frontem

side and three on that side; and the posts thereof and the arches thereof were after the measure of the first gate: the length thereof was fifty cubits, and the breadth five and twenty cubits.

eius et uestibulum eius secundum mensuram portae prioris quinquaginta cubitorum longitudinem eius et latitudinem ui-

40:22 And their windows, and their arches, and their palm ginti quinque cubitorum 40:22 fenestrae autem eius et ue-

trees, were after the measure of the gate that looketh toward the east; and they went up unto it by seven steps; and

the arches thereof were before them. stibulum et scalpturae secundum mensuram portae quae re-

spiciebat ad orientem et septem graduum erat ascensus eius

40:23 And the gate of the inner court was over against the et uestibulum ante eam 40:23 et porta atrii interioris contra

gate toward the north, and toward the east; and he measured from gate to gate an hundred cubits.

portam aquilonis et orientalem et mensus est a porta usque

40:24 After that he brought me toward the south, and be- ad portam centum cubitos 40:24 et duxit me ad uiam auhold a gate toward the south: and he measured the posts thereof and the arches thereof according to these mea-

sures. stralem et ecce porta quae respiciebat ad austrum et mensus

est frontem eius et uestibulum eius iuxta mensuras superio-

40:25 And there were windows in it and in the arches res 40:25 et fenestras eius et uestibula in circuitu sicut fe-

thereof round about, like those windows: the length was fifty cubits, and the breadth five and twenty cubits.

nestras ceteras quinquaginta cubitorum longitudine et latitu-

40:26 And there were seven steps to go up to it, and the dine uiginti quinque cubitorum 40:26 et in gradibus septem

arches thereof were before them: and it had palm trees, one on this side, and another on that side, upon the posts

thereof. ascendebatur ad eam et uestibulum ante fores eius et cela-40:27 And there was a gate in the inner court toward the tae palmae erant una hinc et altera inde in fronte eius 40:27

south: and he measured from gate to gate toward the south an hundred cubits.

et porta atrii interioris in uia australi et mensus est a porta

40:28 And he brought me to the inner court by the south usque ad portam in uia australi centum cubitos 40:28 et in-

gate: and he measured the south gate according to these measures;

Prophetia Ezechielis

troduxit me in atrium interius ad portam australem et mensus est portam iuxta mensuras superiores 40:29 thalamum eius 40:29 And the little chambers thereof, and the posts

thereof, and the arches thereof, according to these measures: and there were windows in it and in the arches thereof round about: it was fifty cubits long, and five and twenty cubits broad.

et frontem eius et uestibulum eius hisdem mensuris et fenestras eius et uestibulorum eius in circuitu quinquaginta cubitos longitudinis et latitudinis uiginti quinque cubitos 40:30 40:30 And the arches round about were five and twenty

cubits long, and five cubits broad.

et uestibulum per gyrum longitudine uiginti quinque cubitorum et latitudine quinque cubitorum 40:31 et uestibulum 40:31 And the arches thereof were toward the utter court;

and palm trees were upon the posts thereof: and the going up to it had eight steps.

eius ad atrium exterius et palmas eius in fronte et octo gradus erant quibus ascendebatur per eam 40:32 et introduxit 40:32 And he brought me into the inner court toward the

east: and he measured the gate according to these measures.

me in atrium interius per uiam orientalem et mensus est portam secundum mensuras superiores 40:33 thalamum eius et 40:33 And the little chambers thereof, and the posts

thereof, and the arches thereof, were according to these measures: and there were windows therein and in the arches thereof round about: it was fifty cubits long, and five and twenty cubits broad.

frontem eius et uestibula eius sicut supra et fenestras eius et uestibuli eius in circuitu longitudine quinquaginta cubitorum et latitudine uiginti quinque cubitorum 40:34 et uestibulum 40:34 And the arches thereof were toward the outward

court; and palm trees were upon the posts thereof, on this side, and on that side: and the going up to it had eight

eius id est atrii exterioris et palmae celatae in fronte eius hinc steps. et inde et in octo gradibus ascensus eius 40:35 et introduxit 40:35 And he brought me to the north gate, and measured

it according to these measures;

me ad portam quae respiciebat ad aquilonem et mensus est

secundum mensuras superiores 40:36 thalamum eius fron- 40:36 The little chambers thereof, the posts thereof, and the arches thereof, and the windows to it round about: the length was fifty cubits, and the breadth five and twenty

tem eius uestibulum eius et fenestras eius per circuitum lon- cubits. gitudine quinquaginta cubitorum et latitudine uiginti quinque cubitorum 40:37 uestibulum eius in atrium exterius et 40:37 And the posts thereof were toward the utter court;

and palm trees were upon the posts thereof, on this side, and on that side: and the going up to it had eight steps.

celatura palmarum in fronte illius hinc et inde et in octo gradibus ascensus eius 40:38 et per singula gazofilacia ostium 40:38 And the chambers and the entries thereof were by

the posts of the gates, where they washed the burnt offering.

in frontibus portarum ibi lauabunt holocaustum 40:39 et in

40:39 And in the porch of the gate were two tables on this side, and two tables on that side, to slay thereon the burnt offering and the sin offering and the trespass offering.

uestibulo portae duae mensae hinc et duae mensae inde ut immoletur super eas holocaustum et pro peccato et pro delicto 40:40 et ad latus exterius quod ascendit ad ostium por- 40:40 And at the side without, as one goeth up to the entry

of the north gate, were two tables; and on the other side, which was at the porch of the gate, were two tables.

tae quae pergit ad aquilonem duae mensae et ad latus alterum ante uestibulum portae duae mensae 40:41 quattuor mensae 40:41 Four tables were on this side, and four tables on that

side, by the side of the gate; eight tables, whereupon they slew their sacrifices.

hinc et quattuor mensae inde per latera portae octo mensae erunt super quas immolabunt 40:42 quattuor autem mensae 40:42 And the four tables were of hewn stone for the burnt

offering, of a cubit and an half long, and a cubit and an half broad, and one cubit high: whereupon also they laid the instruments wherewith they slew the burnt offering and the sacrifice.

ad holocaustum de lapidibus quadris extructae longitudine cubiti unius et dimidii et latitudine cubiti unius et dimidii et

Biblia Sacra Vulgata

altitudine cubiti unius super quas ponant uasa in quibus im-

40:43 And within were hooks, an hand broad, fastened molatur holocaustum et uictima 40:43 et labia earum palmi

round about: and upon the tables was the flesh of the offering.

unius reflexa intrinsecus per circuitum super mensas autem

40:44 And without the inner gate were the chambers of carnes oblationis 40:44 et extra portam interiorem gazofi-

the singers in the inner court, which was at the side of the north gate; and their prospect was toward the south: one at the side of the east gate having the prospect toward the north.

lacia cantorum in atrio interiori quod erat in latere portae respicientis ad aquilonem et facies eorum contra uiam australem una ex latere portae orientalis quae respiciebat ad

40:45 And he said unto me, This chamber, whose prospect uiam aquilonis 40:45 et dixit ad me hoc est gazofilacium is toward the south, is for the priests, the keepers of the charge of the house.

quod respicit uiam meridianam sacerdotum qui excubant in

40:46 And the chamber whose prospect is toward the north custodiis templi 40:46 porro gazofilacium quod respicit ad

is for the priests, the keepers of the charge of the altar: these are the sons of Zadok among the sons of Levi, which

come near to the LORD to minister unto him. uiam aquilonis sacerdotum erit qui excubant ad ministerium

altaris isti sunt filii Sadoc qui accedunt de filiis Leui ad Do-

40:47 So he measured the court, an hundred cubits long, minum ut ministrent ei 40:47 et mensus est atrium longi-

and an hundred cubits broad, foursquare; and the altar that was before the house.

tudine centum cubitorum et latitudine centum cubitorum per

40:48 And he brought me to the porch of the house, and quadrum et altare ante faciem templi 40:48 et introduxit me

measured each post of the porch, five cubits on this side, and five cubits on that side: and the breadth of the gate was three cubits on this side, and three cubits on that side. in uestibulum templi et mensus est uestibulum quinque cu-

bitis hinc et quinque cubitis inde et latitudinem portae trium

40:49 The length of the porch was twenty cubits, and the cubitorum hinc et trium cubitorum inde 40:49 longitudinem

breadth eleven cubits, and he brought me by the steps whereby they went up to it: and there were pillars by the

posts, one on this side, and another on that side. autem uestibuli uiginti cubitorum et latitudinem undecim cubitorum et octo gradibus ascendebatur ad eam et columnae erant in frontibus una hinc et altera inde

41:1 Afterward he brought me to the temple, and mea- 41:1 et introduxit me in templum et mensus est frontes sex

sured the posts, six cubits broad on the one side, and six cubits broad on the other side, which was the breadth of

the tabernacle. cubitos latitudinis hinc et sex cubitos latitudinis inde latitu-41:2 And the breadth of the door was ten cubits; and the dinem tabernaculi 41:2 et latitudo portae decem cubitorum

sides of the door were five cubits on the one side, and five cubits on the other side: and he measured the length

thereof, forty cubits: and the breadth, twenty cubits. erat et latera portae quinque cubitis hinc et quinque cubitis

inde et mensus est longitudinem eius quadraginta cubitorum

41:3 Then went he inward, and measured the post of the et latitudinem uiginti cubitorum 41:3 et introgressus intrin-

door, two cubits; and the door, six cubits; and the breadth of the door, seven cubits.

secus mensus est in fronte portae duos cubitos et portam sex

41:4 So he measured the length thereof, twenty cubits; and cubitorum et latitudinem portae septem cubitorum 41:4 et the breadth, twenty cubits, before the temple: and he said unto me, This is the most holy place.

mensus est longitudinem eius uiginti cubitorum et latitudinem uiginti cubitorum ante faciem templi et dixit ad me hoc

41:5 After he measured the wall of the house, six cubits; est sanctum sanctorum 41:5 et mensus est parietem domus and the breadth of every side chamber, four cubits, round about the house on every side.

sex cubitorum et latitudinem lateris quattuor cubitorum un-

Prophetia Ezechielis

dique per circuitum domus 41:6 latera autem latus ad latus 41:6 And the side chambers were three, one over another,

and thirty in order; and they entered into the wall which was of the house for the side chambers round about, that they might have hold, but they had not hold in the wall of the house.

bis triginta tria et erant eminentia quae ingrederentur per parietem domus in lateribus per circuitum ut continerent et non adtingerent parietem templi 41:7 et platea erat in rotundum 41:7 And there was an enlarging, and a winding about still

upward to the side chambers: for the winding about of the house went still upward round about the house: therefore the breadth of the house was still upward, and so increased from the lowest chamber to the highest by the midst.

ascendens sursum per cocleam et in cenaculum templi deferebat per gyrum idcirco latius erat templum in superioribus et sic de inferioribus ascendebatur ad superiora in medium 41:8 et uidi in domo altitudinem per circuitum fundata la- 41:8 I saw also the height of the house round about: the

foundations of the side chambers were a full reed of six great cubits.

tera ad mensuram calami sex cubitorum spatio 41:9 et la-

41:9 The thickness of the wall, which was for the side chamber without, was five cubits: and that which was left was the place of the side chambers that were within.

titudinem per parietem lateris forinsecus quinque cubitorum et interior domus in lateribus domus 41:10 et inter gazo- 41:10 And between the chambers was the wideness of

twenty cubits round about the house on every side.

filacia latitudinem uiginti cubitorum in circuitu domus undique 41:11 et ostium lateris ad orationem ostium unum ad 41:11 And the doors of the side chambers were toward the

place that was left, one door toward the north, and another door toward the south: and the breadth of the place that

uiam aquilonis et ostium unum ad uiam australem et latitudi- was left was five cubits round about. nem loci ad orationem quinque cubitorum in circuitu 41:12 41:12 Now the building that was before the separate place

at the end toward the west was seventy cubits broad; and the wall of the building was five cubits thick round about,

et aedificium quod erat separatum uersumque ad uiam res- and the length thereof ninety cubits. picientem ad mare latitudinis septuaginta cubitorum paries autem aedificii quinque cubitorum latitudinis per circuitum et longitudo eius nonaginta cubitorum 41:13 et mensus est 41:13 So he measured the house, an hundred cubits long;

and the separate place, and the building, with the walls thereof, an hundred cubits long;

domus longitudinem centum cubitorum et quod separatum erat aedificium et parietes eius longitudinis centum cubito-

rum 41:14 latitudo autem ante faciem domus et eius quod 41:14 Also the breadth of the face of the house, and of the

separate place toward the east, an hundred cubits.

erat separatum contra orientem centum cubitorum 41:15 et 41:15 And he measured the length of the building over

against the separate place which was behind it, and the galleries thereof on the one side and on the other side, an hundred cubits, with the inner temple, and the porches of the court;

mensus est longitudinem aedificii contra faciem eius quod erat separatum ad dorsum ekthetas ex utraque parte centum cubitorum et templum interius et uestibula atrii 41:16 li- 41:16 The door posts, and the narrow windows, and the

galleries round about on their three stories, over against the door, cieled with wood round about, and from the

mina et fenestras obliquas et ekthetas in circuitu per tres par- ground up to the windows, and the windows were covered; tes contra uniuscuiusque limen stratumque ligno per gyrum in circuitu terra autem usque ad fenestras et fenestrae clausae super ostia 41:17 et usque ad domum interiorem et forinse- 41:17 To that above the door, even unto the inner house,

and without, and by all the wall round about within and without, by measure.

cus per omnem parietem in circuitu intrinsecus et forinsecus ad mensuram 41:18 et fabrefacta cherubin et palmae et 41:18 And it was made with cherubims and palm trees, so

that a palm tree was between a cherub and a cherub; and every cherub had two faces;

palma inter cherub et cherub duasque facies habebat cherub

Biblia Sacra Vulgata

41:19 So that the face of a man was toward the palm tree 41:19 faciem hominis iuxta palmam ex hac parte et faciem

on the one side, and the face of a young lion toward the palm tree on the other side: it was made through all the

house round about. leonis iuxta palmam ex alia parte expressam per omnem do-41:20 From the ground unto above the door were cheru- mum in circuitu 41:20 de terra usque ad superiora portae

bims and palm trees made, and on the wall of the temple.

41:21 The posts of the temple were squared, and the face cherubin et palmae celatae erant in pariete templi 41:21 li-

of the sanctuary; the appearance of the one as the appearance of the other.

men quadrangulum et facies sanctuarii aspectus contra as-

41:22 The altar of wood was three cubits high, and the pectum 41:22 altaris lignei trium cubitorum altitudo et lon-

length thereof two cubits; and the corners thereof, and the length thereof, and the walls thereof, were of wood: and

he said unto me, This is the table that is before the LORD. gitudo eius duo cubitorum et anguli eius et longitudo eius et

parietes eius lignei et locutus est ad me haec est mensa co-

41:23 And the temple and the sanctuary had two doors. ram Domino 41:23 et duo ostia erant in templo et in sanc-41:24 And the doors had two leaves apiece, two turning tuario 41:24 et in duobus ostiis ex utraque parte bina erant

leaves; two leaves for the one door, and two leaves for the other door.

ostiola quae in se inuicem plicabantur bina enim ostia erant

41:25 And there were made on them, on the doors of the ex utraque parte ostiorum 41:25 et celata erant in ipsis os-

temple, cherubims and palm trees, like as were made upon the walls; and there were thick planks upon the face of the

porch without. tiis templi cherubin et scalptura palmarum sicut in parietibus

quoque expressa erat quam ob rem erant et grossiora ligna in

41:26 And there were narrow windows and palm trees on uestibuli fronte forinsecus 41:26 super quae fenestrae ob-

the one side and on the other side, on the sides of the porch, and upon the side chambers of the house, and thick

planks. liquae et similitudo palmarum hinc atque inde in umerulis uestibuli secundum latera domus latitudinemque parietum

42:1 Then he brought me forth into the utter court, the way 42:1 et eduxit me in atrium exterius per uiam ducentem toward the north: and he brought me into the chamber that was over against the separate place, and which was before

the building toward the north. ad aquilonem et eduxit me in gazofilacium quod erat contra

separatum aedificium et contra aedem uergentem ad aqui-

42:2 Before the length of an hundred cubits was the north lonem 42:2 in facie longitudinis centum cubitos ostii aqui-

door, and the breadth was fifty cubits.

42:3 Over against the twenty cubits which were for the lonis et latitudinis quinquaginta cubitos 42:3 contra uiginti

inner court, and over against the pavement which was for the utter court, was gallery against gallery in three stories.

cubitos atrii interioris et contra pauimentum stratum lapide

42:4 And before the chambers was a walk to ten cubits atrii exterioris ubi erat porticus iuncta porticui triplici 42:4

breadth inward, a way of one cubit; and their doors toward the north.

et ante gazofilacia deambulatio decem cubitorum latitudinis ad interiora respiciens uiae cubiti unius et ostia earum ad

42:5 Now the upper chambers were shorter: for the gal- aquilonem 42:5 ubi erant gazofilacia in superioribus humi-

leries were higher than these, than the lower, and than the middlemost of the building.

liora quia subportabant porticus quae ex illis eminebant de

42:6 For they were in three stories, but had not pillars as inferioribus et de mediis aedificii 42:6 tristega enim erant et

the pillars of the courts: therefore the building was straitened more than the lowest and the middlemost from the

ground. non habebant columnas sicut erant columnae atriorum prop-42:7 And the wall that was without over against the cham- terea eminebant de inferioribus et de mediis a terra 42:7 et

bers, toward the utter court on the forepart of the chambers, the length thereof was fifty cubits.

peribolus exterior secundum gazofilacia quae erant in uia atrii exterioris ante gazofilacia longitudo eius quinquaginta cu-

Prophetia Ezechielis

bitorum 42:8 quia longitudo erat gazofilaciorum atrii exte- 42:8 For the length of the chambers that were in the utter

court was fifty cubits: and, lo, before the temple were an hundred cubits.

rioris quinquaginta cubitorum et longitudo ante faciem templi centum cubitorum 42:9 et erat subter gazofilacia haec 42:9 And from under these chambers was the entry on the

east side, as one goeth into them from the utter court.

introitus ab oriente ingredientium in ea de atrio exteriori 42:10 in latitudine periboli atrii quod erat contra uiam orien- 42:10 The chambers were in the thickness of the wall of

the court toward the east, over against the separate place, and over against the building.

talem in facie aedificii separati et erant ante aedificium gazofilacia 42:11 et uia ante faciem eorum iuxta similitudi- 42:11 And the way before them was like the appearance

of the chambers which were toward the north, as long as they, and as broad as they: and all their goings out were both according to their fashions, and according to their doors.

nem gazofilaciorum quae erant in uia aquilonis secundum longitudinem eorum sic et latitudo eorum et omnis introitus eorum et similitudines et ostia eorum 42:12 secundum ostia 42:12 And according to the doors of the chambers that

were toward the south was a door in the head of the way, even the way directly before the wall toward the east, as

gazofilaciorum quae erant in uia respiciente ad notum ostium one entereth into them. in capite uiae quae uia erat ante uestibulum separatum per uiam orientalem ingredientibus 42:13 et dixit ad me gazofi- 42:13 Then said he unto me, The north chambers and the

south chambers, which are before the separate place, they be holy chambers, where the priests that approach unto the LORD shall eat the most holy things: there shall they lay the most holy things, and the meat offering, and the sin offering, and the trespass offering; for the place is holy.

lacia aquilonis et gazofilacia austri quae sunt ante aedificium separatum haec sunt gazofilacia sancta in quibus uescuntur sacerdotes qui adpropinquant ad Dominum in sancta sanctorum ibi ponent sancta sanctorum et oblationem pro peccato

et pro delicto locus enim sanctus est 42:14 cum autem in- 42:14 When the priests enter therein, then shall they not go out of the holy place into the utter court, but there they shall lay their garments wherein they minister; for they are holy; and shall put on other garments, and shall approach to those things which are for the people.

gressi fuerint sacerdotes non egredientur de sanctis in atrium exterius et ibi reponent uestimenta sua in quibus ministrant quia sancta sunt uestienturque uestimentis aliis et sic procedent ad populum 42:15 cumque conplesset mensuras do- 42:15 Now when he had made an end of measuring the

inner house, he brought me forth toward the gate whose prospect is toward the east, and measured it round about.

mus interioris eduxit me per uiam portae quae respiciebat ad uiam orientalem et mensus est eam undique per circuitum 42:16 mensus autem est contra uentum orientalem calamo 42:16 He measured the east side with the measuring reed,

five hundred reeds, with the measuring reed round about.

mensurae quingentos calamos in calamo mensurae per circuitum 42:17 et mensus est contra uentum aquilonem quin- 42:17 He measured the north side, five hundred reeds, with

the measuring reed round about.

gentos calamos in calamo mensurae per gyrum 42:18 et ad 42:18 He measured the south side, five hundred reeds,

with the measuring reed.

uentum australem mensus est quingentos calamos in calamo mensurae per circuitum 42:19 et ad uentum occidentalem 42:19 He turned about to the west side, and measured five

hundred reeds with the measuring reed.

mensus est quingentos calamos in calamo mensurae 42:20 42:20 He measured it by the four sides: it had a wall round

about, five hundred reeds long, and five hundred broad, to make a separation between the sanctuary and the profane

per quattuor uentos mensus est illud murum eius undique place. per circuitum longitudine quingentorum cubitorum et latitu-

Biblia Sacra Vulgata

dine quingentorum cubitorum diuidentem inter sanctuarium et uulgi locum

43:1 Afterward he brought me to the gate, even the gate 43:1 et duxit me ad portam quae respiciebat ad uiam orien-

that looketh toward the east:

43:2 And, behold, the glory of the God of Israel came from talem 43:2 et ecce gloria Dei Israhel ingrediebatur per uiam

the way of the east: and his voice was like a noise of many waters: and the earth shined with his glory.

orientalem et uox erat ei quasi uox aquarum multarum et

43:3 And it was according to the appearance of the vision terra splendebat a maiestate eius 43:3 et uidi uisionem se-

which I saw, even according to the vision that I saw when I came to destroy the city: and the visions were like the vision that I saw by the river Chebar; and I fell upon my face.

cundum speciem quam uideram quando uenit ut disperderet ciuitatem et species secundum aspectum quem uideram

43:4 And the glory of the LORD came into the house by iuxta fluuium Chobar et cecidi super faciem meam 43:4 et

the way of the gate whose prospect is toward the east.

maiestas Domini ingressa est templum per uiam portae quae

43:5 So the spirit took me up, and brought me into the respiciebat ad orientem 43:5 et leuauit me spiritus et intro-

inner court; and, behold, the glory of the LORD filled the house.

duxit me in atrium interius et ecce repleta erat gloria Do-

43:6 And I heard him speaking unto me out of the house; mini domus 43:6 et audiui loquentem ad me de domo et and the man stood by me.

43:7 And he said unto me, Son of man, the place of my uir qui stabat iuxta me 43:7 dixit ad me fili hominis locus

throne, and the place of the soles of my feet, where I will dwell in the midst of the children of Israel for ever, and my holy name, shall the house of Israel no more defile, neither they, nor their kings, by their whoredom, nor by the carcases of their kings in their high places.

solii mei et locus uestigiorum pedum meorum ubi habito in medio filiorum Israhel in aeternum et non polluent ultra domus Israhel nomen sanctum meum ipsi et reges eorum in fornicationibus suis et in ruinis regum suorum et in excelsis

43:8 In their setting of their threshold by my thresholds, 43:8 qui fabricati sunt limen suum iuxta limen meum et post-

and their post by my posts, and the wall between me and them, they have even defiled my holy name by their abominations that they have committed: wherefore I have consumed them in mine anger.

es suos iuxta postes meos et murus erat inter me et eos et polluerunt nomen sanctum meum in abominationibus quas

43:9 Now let them put away their whoredom, and the car- fecerunt propter quod consumpsi eos in ira mea 43:9 nunc

cases of their kings, far from me, and I will dwell in the midst of them for ever.

ergo repellant procul fornicationem suam et ruinas regum

43:10 Thou son of man, shew the house to the house of suorum a me et habitabo in medio eorum semper 43:10 tu

Israel, that they may be ashamed of their iniquities: and let them measure the pattern.

autem fili hominis ostende domui Israhel templum et con-

43:11 And if they be ashamed of all that they have done, fundantur ab iniquitatibus suis et metiantur fabricam 43:11

shew them the form of the house, and the fashion thereof, and the goings out thereof, and the comings in thereof, and all the forms thereof, and all the ordinances thereof, and all the forms thereof, and all the laws thereof: and write it in their sight, that they may keep the whole form thereof, and all the ordinances thereof, and do them.

et erubescant ex omnibus quae fecerunt figuram domus et fabricae eius exitus et introitus et omnem descriptionem eius et uniuersa praecepta eius cunctumque ordinem eius et omnes leges eius ostende eis et scribes in oculis eorum et custodiant omnes descriptiones eius et praecepta illius et fa-

43:12 This is the law of the house; Upon the top of the ciant ea 43:12 ista est lex domus in summitate montis om-

mountain the whole limit thereof round about shall be most holy. Behold, this is the law of the house.

nes fines eius in circuitu sanctum sanctorum est haec ergo

Prophetia Ezechielis

est lex domus 43:13 istae autem mensurae altaris in cubito 43:13 And these are the measures of the altar after the cu-

bits: The cubit is a cubit and an hand breadth; even the bottom shall be a cubit, and the breadth a cubit, and the border thereof by the edge thereof round about shall be a span: and this shall be the higher place of the altar.

uerissimo qui habebat cubitum et palmum in sinu eius erat cubitus et cubitus in latitudine et definitio usque ad labium eius in circuitu palmus unus haec quoque erat fossa altaris 43:14 et de sinu terrae usque ad crepidinem nouissimam duo 43:14 And from the bottom upon the ground even to the

lower settle shall be two cubits, and the breadth one cubit; and from the lesser settle even to the greater settle shall be

cubiti et latitudo cubiti unius et a crepidine maiori usque ad four cubits, and the breadth one cubit. crepidinem minorem quattuor cubiti et latitudo unius cubiti 43:15 ipse autem arihel quattuor cubitorum et ab arihel us- 43:15 So the altar shall be four cubits; and from the altar

and upward shall be four horns.

que sursum cornua quattuor 43:16 et arihel duodecim cubi- 43:16 And the altar shall be twelve cubits long, twelve

broad, square in the four squares thereof.

torum in longitudine per duodecim cubitos latitudinis quadrangulatum aequis lateribus 43:17 et crepido quattuorde- 43:17 And the settle shall be fourteen cubits long and

fourteen broad in the four squares thereof; and the border about it shall be half a cubit; and the bottom thereof shall be a cubit about; and his stairs shall look toward the east.

cim cubitorum longitudinis per quattuordecim latitudinis in quattuor angulis eius et corona in circuitu eius dimidii cubitus et sinus eius unius cubiti per circuitum gradus autem eius uersi ad orientem 43:18 et dixit ad me fili hominis haec dicit 43:18 And he said unto me, Son of man, thus saith the

Lord GOD; These are the ordinances of the altar in the day when they shall make it, to offer burnt offerings thereon,

Dominus Deus hii sunt ritus altaris in quacumque die fuerit and to sprinkle blood thereon. fabricatum ut offeratur super illud holocaustum et effundatur sanguis 43:19 et dabis sacerdotibus Leuitis qui sunt de 43:19 And thou shalt give to the priests the Levites that be

of the seed of Zadok, which approach unto me, to minister unto me, saith the Lord GOD, a young bullock for a sin

semine Sadoc qui accedunt ad me ait Dominus Deus ut of- offering. ferant mihi uitulum de armento pro peccato 43:20 et adsu- 43:20 And thou shalt take of the blood thereof, and put it

on the four horns of it, and on the four corners of the settle, and upon the border round about: thus shalt thou cleanse

mens de sanguine eius pones super quattuor cornua eius et and purge it. super quattuor angulos crepidinis et super coronam in circuitu et mundabis illud et expiabis 43:21 et tolles uitulum 43:21 Thou shalt take the bullock also of the sin offering,

and he shall burn it in the appointed place of the house, without the sanctuary.

qui oblatus fuerit pro peccato et conbures illum in separato loco domus extra sanctuarium 43:22 et in die secunda offe- 43:22 And on the second day thou shalt offer a kid of the

goats without blemish for a sin offering; and they shall cleanse the altar, as they did cleanse it with the bullock.

res hircum caprarum inmaculatum pro peccato et expiabunt altare sicut expiauerunt in uitulo 43:23 cumque conpleue- 43:23 When thou hast made an end of cleansing it, thou

shalt offer a young bullock without blemish, and a ram out of the flock without blemish.

ris expians illud offeres uitulum de armento inmaculatum et arietem de grege inmaculatum 43:24 et offeres eos in con- 43:24 And thou shalt offer them before the LORD, and the

priests shall cast salt upon them, and they shall offer them up for a burnt offering unto the LORD.

spectu Domini et mittent sacerdotes super eos sal et offerent eos holocaustum Domino 43:25 septem diebus facies hir- 43:25 Seven days shalt thou prepare every day a goat for a

sin offering: they shall also prepare a young bullock, and a ram out of the flock, without blemish.

cum pro peccato cotidie et uitulum de armento et arietem de pecoribus inmaculatos offerent 43:26 septem diebus ex- 43:26 Seven days shall they purge the altar and purify it;

and they shall consecrate themselves.

Biblia Sacra Vulgata

piabunt altare et mundabunt illud et implebunt manum eius

43:27 And when these days are expired, it shall be, that 43:27 expletis autem diebus in die octaua et ultra facient sa-

upon the eighth day, and so forward, the priests shall make your burnt offerings upon the altar, and your peace offer-

ings; and I will accept you, saith the Lord GOD. cerdotes super altare holocausta uestra et quae pro pace offerunt et placatus ero uobis ait Dominus Deus

44:1 Then he brought me back the way of the gate of the 44:1 et conuertit me ad uiam portae sanctuarii exterioris

outward sanctuary which looketh toward the east; and it was shut.

quae respiciebat ad orientem et erat clausa 44:2 et dixit Do-

44:2 Then said the LORD unto me; This gate shall be shut, it shall not be opened, and no man shall enter in by it; because the LORD, the God of Israel, hath entered in by it, therefore it shall be shut.

minus ad me porta haec clausa erit non aperietur et uir non transiet per eam quoniam Dominus Deus Israhel ingressus

44:3 It is for the prince; the prince, he shall sit in it to eat est per eam eritque clausa 44:3 principi princeps ipse sede-

bread before the LORD; he shall enter by the way of the porch of that gate, and shall go out by the way of the same.

bit in ea ut comedat panem coram Domino per uiam uesti-

44:4 Then brought he me the way of the north gate be- buli portae ingredietur et per uiam eius egredietur 44:4 et

fore the house: and I looked, and, behold, the glory of the LORD filled the house of the LORD: and I fell upon my

face. adduxit me per uiam portae aquilonis in conspectu domus et

uidi et ecce impleuit gloria Domini domum Domini et ce-

44:5 And the LORD said unto me, Son of man, mark well, cidi in faciem meam 44:5 et dixit ad me Dominus fili ho-

and behold with thine eyes, and hear with thine ears all that I say unto thee concerning all the ordinances of the house of the LORD, and all the laws thereof; and mark well the entering in of the house, with every going forth of the sanctuary.

minis pone cor tuum et uide oculis tuis et auribus tuis audi omnia quae ego loquor ad te de uniuersis caerimoniis domus Domini et de cunctis legibus eius et pones cor tuum

44:6 And thou shalt say to the rebellious, even to the house in uiis templi per omnes exitus sanctuarii 44:6 et dices ad of Israel, Thus saith the Lord GOD; O ye house of Israel, let it suffice you of all your abominations,

exasperantem me domum Israhel haec dicit Dominus De-

44:7 In that ye have brought into my sanctuary strangers, us sufficiant uobis omnia scelera uestra domus Israhel 44:7

uncircumcised in heart, and uncircumcised in flesh, to be in my sanctuary, to pollute it, even my house, when ye offer my bread, the fat and the blood, and they have broken my covenant because of all your abominations.

eo quod inducitis filios alienos incircumcisos corde et incircumcisos carne ut sint in sanctuario meo et polluant domum meam et offertis panes meos adipem et sanguinem et dis-

44:8 And ye have not kept the charge of mine holy things: soluitis pactum meum in omnibus sceleribus uestris 44:8

but ye have set keepers of my charge in my sanctuary for yourselves.

et non seruastis praecepta sanctuarii mei et posuistis custodes obseruationum mearum in sanctuario meo uobismet

44:9 Thus saith the Lord GOD; No stranger, uncircum- ipsis 44:9 haec dicit Dominus Deus omnis alienigena incir-

cised in heart, nor uncircumcised in flesh, shall enter into my sanctuary, of any stranger that is among the children

of Israel. cumcisus corde et incircumcisus carne non ingredietur sanc-

tuarium meum omnis filius alienus qui est in medio filio-

44:10 And the Levites that are gone away far from me, rum Israhel 44:10 sed et Leuitae qui longe recesserunt a me

when Israel went astray, which went astray away from me after their idols; they shall even bear their iniquity.

in errore filiorum Israhel et errauerunt a me post idola sua

44:11 Yet they shall be ministers in my sanctuary, having et portauerunt iniquitatem suam 44:11 erunt in sanctuario

charge at the gates of the house, and ministering to the house: they shall slay the burnt offering and the sacrifice for the people, and they shall stand before them to minister unto them.

meo aeditui et ianitores portarum domus et ministri domus

Prophetia Ezechielis

ipsi mactabunt holocaustosin et uictimas populi et ipsi stabunt in conspectu eorum ut ministrent eis 44:12 pro eo quod 44:12 Because they ministered unto them before their

idols, and caused the house of Israel to fall into iniquity; therefore have I lifted up mine hand against them, saith the

ministrauerunt illis in conspectu idolorum suorum et facti Lord GOD, and they shall bear their iniquity. sunt domui Israhel in offendiculum iniquitatis idcirco leuaui manum meam super eos dicit Dominus Deus et portauerunt iniquitatem suam 44:13 et non adpropinquabunt ad me ut 44:13 And they shall not come near unto me, to do the

office of a priest unto me, nor to come near to any of my holy things, in the most holy place: but they shall bear their shame, and their abominations which they have committed.

sacerdotio fungantur mihi neque accedent ad omne sanctuarium meum iuxta sancta sanctorum sed portabunt confusionem suam et scelera sua quae fecerunt 44:14 et dabo eos 44:14 But I will make them keepers of the charge of the

house, for all the service thereof, and for all that shall be done therein.

ianitores domus in omni ministerio eius et uniuersis quae

fiunt in ea 44:15 sacerdotes autem Leuitae filii Sadoc qui 44:15 But the priests the Levites, the sons of Zadok, that kept the charge of my sanctuary when the children of Israel went astray from me, they shall come near to me to minister unto me, and they shall stand before me to offer unto me the fat and the blood, saith the Lord GOD:

custodierunt caerimonias sanctuarii mei cum errarent filii Israhel a me ipsi accedent ad me ut ministrent mihi et stabunt in conspectu meo ut offerant mihi adipem et sanguinem ait Dominus Deus 44:16 ipsi ingredientur sanctuarium meum 44:16 They shall enter into my sanctuary, and they shall

come near to my table, to minister unto me, and they shall keep my charge.

et ipsi accedent ad mensam meam ut ministrent mihi et custodiant caerimonias meas 44:17 cumque ingredientur por- 44:17 And it shall come to pass, that when they enter in

at the gates of the inner court, they shall be clothed with linen garments; and no wool shall come upon them, whiles

tas atrii interioris uestibus lineis induentur nec ascendet su- they minister in the gates of the inner court, and within. per eos quicquam laneum quando ministrant in portis atrii interioris et intrinsecus 44:18 uittae lineae erunt in capiti- 44:18 They shall have linen bonnets upon their heads, and

shall have linen breeches upon their loins; they shall not gird themselves with any thing that causeth sweat.

bus eorum et feminalia linea erunt in lumbis eorum et non accingentur in sudore 44:19 cumque egredientur atrium ex- 44:19 And when they go forth into the utter court, even

into the utter court to the people, they shall put off their garments wherein they ministered, and lay them in the holy chambers, and they shall put on other garments; and they shall not sanctify the people with their garments.

terius ad populum exuent se uestimenta sua in quibus ministrauerunt et reponent ea in gazofilacio sanctuarii et uestient se uestimentis aliis et non sanctificabunt populum in uesti-

bus suis 44:20 caput autem suum non radent neque comam 44:20 Neither shall they shave their heads, nor suffer their

locks to grow long; they shall only poll their heads.

nutrient sed tondentes adtondent capita sua 44:21 et uinum 44:21 Neither shall any priest drink wine, when they enter

into the inner court.

non bibet omnis sacerdos quando ingressurus est atrium interius 44:22 et uiduam et repudiatam non accipient uxores 44:22 Neither shall they take for their wives a widow, nor

her that is put away: but they shall take maidens of the seed of the house of Israel, or a widow that had a priest

sed uirgines de semine domus Israhel sed et uiduam quae before. fuerit uidua a sacerdote accipient 44:23 et populum meum 44:23 And they shall teach my people the difference be-

tween the holy and profane, and cause them to discern between the unclean and the clean.

docebunt quid sit inter sanctum et pollutum et inter mun-

dum et inmundum ostendent eis 44:24 et cum fuerit con- 44:24 And in controversy they shall stand in judgment; and they shall judge it according to my judgments: and they shall keep my laws and my statutes in all mine assemblies; and they shall hallow my sabbaths.

Biblia Sacra Vulgata

trouersia stabunt in iudiciis meis et iudicabunt leges meas et praecepta mea in omnibus sollemnitatibus meis custodient et

44:25 And they shall come at no dead person to defile sabbata mea sanctificabunt 44:25 et ad mortuum hominem

themselves: but for father, or for mother, or for son, or for daughter, for brother, or for sister that hath had no hus-

band, they may defile themselves. non ingredientur ne polluantur nisi ad patrem et matrem et

filium et filiam et fratrem et sororem quae alterum uirum

44:26 And after he is cleansed, they shall reckon unto him non habuit in quibus contaminabuntur 44:26 et postquam seven days. 44:27 And in the day that he goeth into the sanctuary, unto fuerit emundatus septem dies numerabuntur ei 44:27 et in the inner court, to minister in the sanctuary, he shall offer his sin offering, saith the Lord GOD.

die introitus sui in sanctuarium ad atrium interius ut ministret mihi in sanctuario offeret pro peccato suo ait Dominus

44:28 And it shall be unto them for an inheritance: I am Deus 44:28 erit autem eis hereditas ego hereditas eorum et

their inheritance: and ye shall give them no possession in Israel: I am their possession.

possessionem non dabitis eis in Israhel ego enim possessio

44:29 They shall eat the meat offering, and the sin offer- eorum 44:29 uictimam et pro peccato et pro delicto ipsi co-

ing, and the trespass offering: and every dedicated thing in Israel shall be theirs.

medent et omne uotum in Israhel ipsorum erit 44:30 et pri-

44:30 And the first of all the firstfruits of all things, and every oblation of all, of every sort of your oblations, shall be the priest’s: ye shall also give unto the priest the first of your dough, that he may cause the blessing to rest in thine house.

mitiua omnium primogenitorum et omnia libamenta ex omnibus quae offeruntur sacerdotum erunt et primitiua ciborum uestrorum dabitis sacerdoti ut reponat benedictionem domui

44:31 The priests shall not eat of any thing that is dead of suae 44:31 omne morticinum et captum a bestia de auibus itself, or torn, whether it be fowl or beast.

et de pecoribus non comedent sacerdotes

45:1 Moreover, when ye shall divide by lot the land for 45:1 cumque coeperitis terram diuidere sortito separate

inheritance, ye shall offer an oblation unto the LORD, an holy portion of the land: the length shall be the length of five and twenty thousand reeds, and the breadth shall be ten thousand. This shall be holy in all the borders thereof round about.

primitias Domino sanctificatum de terra longitudine uiginti quinque milia et latitudine decem milia sanctificatum erit

45:2 Of this there shall be for the sanctuary five hundred in in omni termino eius per circuitum 45:2 et erit ex omni length, with five hundred in breadth, square round about; and fifty cubits round about for the suburbs thereof.

parte sanctificatum quingentos per quingentos quadrifariam per circuitum et quinquaginta cubitis in suburbana eius per

45:3 And of this measure shalt thou measure the length of gyrum 45:3 et a mensura ista mensurabis longitudinem ui-

five and twenty thousand, and the breadth of ten thousand: and in it shall be the sanctuary and the most holy place.

ginti quinque milium et latitudinem decem milium et in ipso

45:4 The holy portion of the land shall be for the priests erit templum sanctumque sanctorum 45:4 sanctificatum de

the ministers of the sanctuary, which shall come near to minister unto the LORD: and it shall be a place for their

houses, and an holy place for the sanctuary. terra erit sacerdotibus ministris sanctuarii qui accedunt ad

ministerium Domini et erit eis locus in domos et in sanctua-

45:5 And the five and twenty thousand of length, and the rium sanctitatis 45:5 uiginti quinque autem milia longitu-

ten thousand of breadth shall also the Levites, the ministers of the house, have for themselves, for a possession for

twenty chambers. dinis et decem milia latitudinis erunt Leuitis qui ministrant 45:6 And ye shall appoint the possession of the city five domui ipsi possidebunt uiginti gazofilacia 45:6 et posses-

thousand broad, and five and twenty thousand long, over against the oblation of the holy portion: it shall be for the

whole house of Israel. sionem ciuitatis dabitis quinque milia latitudinis et longitu-

Prophetia Ezechielis

dinis uiginti quinque milia secundum separationem sanctuarii omni domui Israhel 45:7 principi quoque hinc et inde in 45:7 And a portion shall be for the prince on the one side

and on the other side of the oblation of the holy portion, and of the possession of the city, before the oblation of the holy portion, and before the possession of the city, from the west side westward, and from the east side eastward: and the length shall be over against one of the portions, from the west border unto the east border.

separationem sanctuarii et in possessionem ciuitatis contra faciem separationis sanctuarii et contra faciem possessionis urbis a latere maris usque ad mare et a latere orientis usque ad orientem longitudinem autem iuxta unamquamque partium a termino occidentali usque ad terminum orientalem 45:8 de terra erit ei possessio in Israhel et non depopulabun- 45:8 In the land shall be his possession in Israel: and my

princes shall no more oppress my people; and the rest of the land shall they give to the house of Israel according to

tur ultra principes populum meum sed terram dabunt domui their tribes. Israhel secundum tribus eorum 45:9 haec dicit Dominus 45:9 Thus saith the Lord GOD; Let it suffice you, O

princes of Israel: remove violence and spoil, and execute judgment and justice, take away your exactions from my

Deus sufficiat uobis principes Israhel iniquitatem et rapinas people, saith the Lord GOD. intermittite et iudicium et iustitiam facite separate confinia uestra a populo meo ait Dominus Deus 45:10 statera iusta et 45:10 Ye shall have just balances, and a just ephah, and a

just bath.

oephi iustum et batus iustus erit uobis 45:11 oephi et batus 45:11 The ephah and the bath shall be of one measure,

that the bath may contain the tenth part of an homer, and the ephah the tenth part of an homer: the measure thereof

aequalia et unius mensurae erunt ut capiat decimam partem shall be after the homer. chori batus et decimam partem chori oephi iuxta mensuram chori erit aequa libratio eorum 45:12 siclus autem uiginti 45:12 And the shekel shall be twenty gerahs: twenty

shekels, five and twenty shekels, fifteen shekels, shall be your maneh.

obolos habeat porro uiginti sicli et uiginti quinque sicli et quindecim sicli minam facient 45:13 et haec sunt primitiae 45:13 This is the oblation that ye shall offer; the sixth part

of an ephah of an homer of wheat, and ye shall give the sixth part of an ephah of an homer of barley:

quas tolletis sextam partem oephi de choro frumenti et sex-

tam partem oephi de choro hordei 45:14 mensura quoque 45:14 Concerning the ordinance of oil, the bath of oil, ye shall offer the tenth part of a bath out of the cor, which is an homer of ten baths; for ten baths are an homer:

olei batus olei decima pars chori est et decem bati chorum faciunt quia decem bati implent chorum 45:15 et arietem 45:15 And one lamb out of the flock, out of two hundred,

out of the fat pastures of Israel; for a meat offering, and for a burnt offering, and for peace offerings, to make rec-

unum de grege ducentorum de his quae nutriunt Israhel in onciliation for them, saith the Lord GOD. sacrificium et in holocaustum et in pacifica ad expiandum

pro eis ait Dominus Deus 45:16 omnis populus terrae tene- 45:16 All the people of the land shall give this oblation for

the prince in Israel.

bitur primitiis his principi in Israhel 45:17 et super princi- 45:17 And it shall be the prince’s part to give burnt offer-

ings, and meat offerings, and drink offerings, in the feasts, and in the new moons, and in the sabbaths, in all solemnities of the house of Israel: he shall prepare the sin offering, and the meat offering, and the burnt offering, and the peace offerings, to make reconciliation for the house of Israel.

pem erunt holocausta et sacrificium et libamina in sollemnitatibus et in kalendis et in sabbatis in uniuersis sollemnitatibus domus Israhel ipse faciat pro peccato sacrificium et holocaustum et pacifica ad expiandum pro domo Israhel 45:18 haec dicit Dominus Deus in primo mense una mensis 45:18 Thus saith the Lord GOD; In the first month, in the

first day of the month, thou shalt take a young bullock without blemish, and cleanse the sanctuary:

sumes uitulum de armento inmaculatum et expiabis sanc-

Biblia Sacra Vulgata

45:19 And the priest shall take of the blood of the sin of- tuarium 45:19 et tollet sacerdos de sanguine quod erit pro

fering, and put it upon the posts of the house, and upon the four corners of the settle of the altar, and upon the posts of

the gate of the inner court. peccato et ponet in postibus domus et in quattuor angulis 45:20 And so thou shalt do the seventh day of the month crepidinis altaris et in postibus portae atrii interioris 45:20

for every one that erreth, and for him that is simple: so shall ye reconcile the house.

et sic facies in septima mensis pro unoquoque qui ignorauit

45:21 In the first month, in the fourteenth day of the et errore deceptus est et expiabitis pro domo 45:21 in primo

month, ye shall have the passover, a feast of seven days; unleavened bread shall be eaten.

mense quartadecima die mensis erit uobis paschae sollemni-

45:22 And upon that day shall the prince prepare for him- tas septem diebus azyma comedentur 45:22 et faciet prin-

self and for all the people of the land a bullock for a sin offering.

ceps in die illa pro se et pro uniuerso populo terrae uitulum

45:23 And seven days of the feast he shall prepare a burnt pro peccato 45:23 et in septem dierum sollemnitate faciet

offering to the LORD, seven bullocks and seven rams without blemish daily the seven days; and a kid of the

goats daily for a sin offering. holocaustum Domino septem uitulos et septem arietes inma-

culatos cotidie septem diebus et pro peccato hircum capra-

45:24 And he shall prepare a meat offering of an ephah for rum cotidie 45:24 et sacrificium oephi per uitulum et oephi

a bullock, and an ephah for a ram, and an hin of oil for an ephah.

per arietem faciet et olei hin per singula oephi 45:25 sep-

45:25 In the seventh month, in the fifteenth day of the month, shall he do the like in the feast of the seven days, according to the sin offering, according to the burnt offering, and according to the meat offering, and according to the oil.

timo mense quintadecima die mensis in sollemnitate faciet sicut supra dicta sunt per septem dies tam pro peccato quam pro holocausto et in sacrificio et in oleo

46:1 Thus saith the Lord GOD; The gate of the inner court 46:1 haec dicit Dominus Deus porta atrii interioris quae re-

that looketh toward the east shall be shut the six working days; but on the sabbath it shall be opened, and in the day

of the new moon it shall be opened. spicit ad orientem erit clausa sex diebus in quibus opus fit die

autem sabbati aperietur sed et in die kalendarum aperietur

46:2 And the prince shall enter by the way of the porch 46:2 et intrabit princeps per uiam uestibuli portae de foris et

of that gate without, and shall stand by the post of the gate, and the priests shall prepare his burnt offering and his peace offerings, and he shall worship at the threshold of the gate: then he shall go forth; but the gate shall not be shut until the evening.

stabit in limine portae et facient sacerdotes holocaustum eius et pacifica eius et adorabit super limen portae et egredietur

46:3 Likewise the people of the land shall worship at the porta autem non claudetur usque ad uesperam 46:3 et addoor of this gate before the LORD in the sabbaths and in the new moons.

orabit populus terrae ad ostium portae illius in sabbatis et in

46:4 And the burnt offering that the prince shall offer unto kalendis coram Domino 46:4 holocaustum autem hoc offethe LORD in the sabbath day shall be six lambs without blemish, and a ram without blemish.

ret princeps Domino in die sabbati sex agnos inmaculatos et

46:5 And the meat offering shall be an ephah for a ram, arietem inmaculatum 46:5 et sacrificium oephi per arietem

and the meat offering for the lambs as he shall be able to give, and an hin of oil to an ephah.

agnis autem sacrificium quod dederit manus eius et olei hin

46:6 And in the day of the new moon it shall be a young per singula oephi 46:6 in die autem kalendarum uitulum de

bullock without blemish, and six lambs, and a ram: they shall be without blemish.

armento inmaculatum et sex agni et arietes inmaculati erunt

46:7 And he shall prepare a meat offering, an ephah for a 46:7 et oephi per uitulum oephi quoque per arietem faciet sa-

bullock, and an ephah for a ram, and for the lambs according as his hand shall attain unto, and an hin of oil to an

ephah. crificium agnis autem sicut inuenerit manus eius et olei hin 46:8 And when the prince shall enter, he shall go in by the per singula oephi 46:8 cumque ingressurus est princeps per

way of the porch of that gate, and he shall go forth by the way thereof.

Prophetia Ezechielis

uiam uestibuli portae ingrediatur et per eandem uiam exeat 46:9 et cum intrabit populus terrae in conspectu Domini in 46:9 But when the people of the land shall come before

the LORD in the solemn feasts, he that entereth in by the way of the north gate to worship shall go out by the way of the south gate; and he that entereth by the way of the south gate shall go forth by the way of the north gate: he shall not return by the way of the gate whereby he came in, but shall go forth over against it.

sollemnitatibus qui ingreditur per portam aquilonis ut adoret egrediatur per uiam portae meridianae porro qui ingreditur per uiam portae meridianae egrediatur per uiam portae aquilonis non reuertetur per uiam portae per quam ingressus est sed e regione illius egredietur 46:10 princeps autem in me- 46:10 And the prince in the midst of them, when they go

in, shall go in; and when they go forth, shall go forth.

dio eorum cum ingredientibus ingredietur et cum egredientibus egredietur 46:11 et in nundinis et in sollemnitatibus 46:11 And in the feasts and in the solemnities the meat

offering shall be an ephah to a bullock, and an ephah to a ram, and to the lambs as he is able to give, and an hin of

erit sacrificium oephi per uitulum et oephi per arietem agnis oil to an ephah. autem erit sacrificium sicut inuenerit manus eius et olei hin per singula oephi 46:12 cum autem fecerit princeps spon- 46:12 Now when the prince shall prepare a voluntary burnt

offering or peace offerings voluntarily unto the LORD, one shall then open him the gate that looketh toward the east, and he shall prepare his burnt offering and his peace offerings, as he did on the sabbath day: then he shall go forth; and after his going forth one shall shut the gate.

taneum holocaustum aut pacifica uoluntaria Domino aperietur ei porta quae respicit ad orientem et faciet holocaustum suum et pacifica sua sicut fieri solet in die sabbati et egredietur claudeturque porta postquam exierit 46:13 et agnum 46:13 Thou shalt daily prepare a burnt offering unto the

LORD of a lamb of the first year without blemish: thou shalt prepare it every morning.

eiusdem anni inmaculatum faciet holocaustum cotidie Domino semper mane faciet illud 46:14 et sacrificium faciet 46:14 And thou shalt prepare a meat offering for it every

morning, the sixth part of an ephah, and the third part of an hin of oil, to temper with the fine flour; a meat offering

super eo cata mane mane sextam partem oephi et de oleo continually by a perpetual ordinance unto the LORD. tertiam partem hin ut misceatur similae sacrificium Domino legitimum iuge atque perpetuum 46:15 faciet agnum et sa- 46:15 Thus shall they prepare the lamb, and the meat of-

fering, and the oil, every morning for a continual burnt offering.

crificium et oleum cata mane mane holocaustum sempiternum 46:16 haec dicit Dominus Deus si dederit princeps do- 46:16 Thus saith the Lord GOD; If the prince give a gift

unto any of his sons, the inheritance thereof shall be his sons’; it shall be their possession by inheritance.

num alicui de filiis suis hereditas eius filiorum suorum erit possidebunt ea hereditarie 46:17 si autem dederit legatum 46:17 But if he give a gift of his inheritance to one of his

servants, then it shall be his to the year of liberty; after it shall return to the prince: but his inheritance shall be his

de hereditate sua uni seruorum suorum erit illius usque ad sons’ for them. annum remissionis et reuertetur ad principem hereditas autem eius filiis eius erit 46:18 et non accipiet princeps de he- 46:18 Moreover the prince shall not take of the people’s

inheritance by oppression, to thrust them out of their possession; but he shall give his sons inheritance out of his own possession: that my people be not scattered every man from his possession.

reditate populi per uiolentiam et de possessione eorum sed de possessione sua hereditatem dabit filiis suis ut non dispergatur populus meus unusquisque a possessione sua 46:19 et 46:19 After he brought me through the entry, which was at

the side of the gate, into the holy chambers of the priests, which looked toward the north: and, behold, there was a

introduxit me per ingressum qui erat ex latere portae in ga- place on the two sides westward. zofilacia sanctuarii ad sacerdotes quae respiciebant ad aqui-

Biblia Sacra Vulgata

46:20 Then said he unto me, This is the place where the lonem et erat ibi locus uergens ad occidentem 46:20 et dixit

priests shall boil the trespass offering and the sin offering, where they shall bake the meat offering; that they bear

them not out into the utter court, to sanctify the people. ad me iste est locus ubi coquent sacerdotes pro delicto et pro

peccato ubi coquent sacrificium ut non efferant in atrio ex-

46:21 Then he brought me forth into the utter court, and teriori et sanctificetur populus 46:21 et eduxit me in atrium

caused me to pass by the four corners of the court; and, behold, in every corner of the court there was a court.

exterius et circumduxit me per quattuor angulos atrii et ecce atriolum erat in angulo atrii atriola singula per angulos atrii

46:22 In the four corners of the court there were courts 46:22 in quattuor angulos atrii atriola disposita quadraginta

joined of forty cubits long and thirty broad: these four corners were of one measure.

cubitorum per longum et triginta per latum mensurae unius

46:23 And there was a row of building round about in quattuor erant 46:23 et paries per circuitum ambiens quat-

them, round about them four, and it was made with boiling places under the rows round about.

tuor atriola et culinae fabricatae erant subter porticus per gy-

46:24 Then said he unto me, These are the places of them rum 46:24 et dixit ad me haec est domus culinarum in qua that boil, where the ministers of the house shall boil the sacrifice of the people.

coquent ministri domus Domini uictimas populi

47:1 Afterward he brought me again unto the door of 47:1 et conuertit me ad portam domus et ecce aquae egre-

the house; and, behold, waters issued out from under the threshold of the house eastward: for the forefront of the house stood toward the east, and the waters came down from under from the right side of the house, at the south side of the altar.

diebantur subter limen domus ad orientem facies enim domus respiciebat ad orientem aquae autem descendebant in

47:2 Then brought he me out of the way of the gate north- latus templi dextrum ad meridiem altaris 47:2 et eduxit me

ward, and led me about the way without unto the utter gate by the way that looketh eastward; and, behold, there ran

out waters on the right side. per uiam portae aquilonis et conuertit me ad uiam foras por-

tam exteriorem uiam quae respiciebat ad orientem et ecce

47:3 And when the man that had the line in his hand went aquae redundantes a latere dextro 47:3 cum egrederetur uir

forth eastward, he measured a thousand cubits, and he brought me through the waters; the waters were to the an-

kles. ad orientem qui habebat funiculum in manu sua et mensus

est mille cubitos et transduxit me per aquam usque ad talos

47:4 Again he measured a thousand, and brought me 47:4 rursumque mensus est mille et transduxit me per aquam

through the waters; the waters were to the knees. Again he measured a thousand, and brought me through; the wa-

ters were to the loins. usque ad genua 47:5 et mensus est mille et transduxit me

47:5 Afterward he measured a thousand; and it was a river that I could not pass over: for the waters were risen, waters to swim in, a river that could not be passed over.

per aquam usque ad renes et mensus est mille torrentem quem non potui pertransire quoniam intumuerant aquae pro-

47:6 And he said unto me, Son of man, hast thou seen this? fundae torrentis qui non potest transuadari 47:6 et dixit ad

Then he brought me, and caused me to return to the brink of the river.

me certe uidisti fili hominis et duxit me et conuertit ad ri-

47:7 Now when I had returned, behold, at the bank of the pam torrentis 47:7 cumque me conuertissem ecce in ripa

river were very many trees on the one side and on the other.

47:8 Then said he unto me, These waters issue out toward torrentis ligna multa nimis ex utraque parte 47:8 et ait ad

the east country, and go down into the desert, and go into the sea: which being brought forth into the sea, the waters

shall be healed. me aquae istae quae egrediuntur ad tumulos sabuli orientalis

et descendunt ad plana deserti intrabunt mare et exibunt et

47:9 And it shall come to pass, that every thing that liveth, sanabuntur aquae 47:9 et omnis anima uiuens quae serpit which moveth, whithersoever the rivers shall come, shall live: and there shall be a very great multitude of fish, because these waters shall come thither: for they shall be healed; and every thing shall live whither the river cometh.

quocumque uenerit torrens uiuet et erunt pisces multi satis

Prophetia Ezechielis

postquam uenerint illuc aquae istae et sanabuntur et uiuent

omnia ad quae uenerit torrens 47:10 uiuent et stabunt su- 47:10 And it shall come to pass, that the fishers shall stand upon it from Engedi even unto Eneglaim; they shall be a place to spread forth nets; their fish shall be according to

per illa piscatores ab Engaddi usque ad Engallim siccatio their kinds, as the fish of the great sea, exceeding many. sagenarum erunt plurimae species erunt piscium eius sicut pisces maris magni multitudinis nimiae 47:11 in litoribus 47:11 But the miry places thereof and the marishes thereof

shall not be healed; they shall be given to salt.

autem eius et in palustribus non sanabuntur quia in salinas dabuntur 47:12 et super torrentem orietur in ripis eius ex 47:12 And by the river upon the bank thereof, on this side

and on that side, shall grow all trees for meat, whose leaf shall not fade, neither shall the fruit thereof be consumed: it shall bring forth new fruit according to his months, because their waters they issued out of the sanctuary: and the fruit thereof shall be for meat, and the leaf thereof for medicine.

utraque parte omne lignum pomiferum non defluet folium ex eo et non deficiet fructus eius per singulos menses adferet primitiua quia aquae eius de sanctuario egredientur et erunt fructus eius in cibum et folia eius ad medicinam 47:13 haec 47:13 Thus saith the Lord GOD; This shall be the border,

whereby ye shall inherit the land according to the twelve tribes of Israel: Joseph shall have two portions.

dicit Dominus Deus hic est terminus in quo possidebitis terram in duodecim tribubus Israhel quia Ioseph duplicem funiculum habet 47:14 possidebitis autem eam singuli aeque 47:14 And ye shall inherit it, one as well as another: con-

cerning the which I lifted up mine hand to give it unto your fathers: and this land shall fall unto you for inheritance.

ut frater suus quam leuaui manum meam ut darem patribus uestris et cadet terra haec uobis in possessionem 47:15 hic 47:15 And this shall be the border of the land toward the

north side, from the great sea, the way of Hethlon, as men go to Zedad;

est autem terminus terrae ad plagam septentrionalem a mari magno uia Bethalon uenientibus Sadada 47:16 Emath Be- 47:16 Hamath, Berothah, Sibraim, which is between the

border of Damascus and the border of Hamath; Hazarhatticon, which is by the coast of Hauran.

rotha Sabarim quae est inter terminum Damasci et confinium Emath domus Atticon quae est iuxta terminos Auran 47:17 47:17 And the border from the sea shall be Hazarenan,

the border of Damascus, and the north northward, and the border of Hamath. And this is the north side.

et erit terminus a mari usque ad atrium Aenon terminus Damasci et ab aquilone ad aquilonem et terminus Emath plaga autem septentrionalis 47:18 porro plaga orientalis de me- 47:18 And the east side ye shall measure from Hauran,

and from Damascus, and from Gilead, and from the land of Israel by Jordan, from the border unto the east sea. And

dio Auran et de medio Damasci et de medio Galaad et de this is the east side. medio terrae Israhel Iordanis disterminans ad mare orientale metiemini etiam plagam orientalem 47:19 plaga autem au- 47:19 And the south side southward, from Tamar even to

the waters of strife in Kadesh, the river to the great sea. And this is the south side southward.

stralis meridiana a Thamar usque ad aquas Contradictionis Cades et torrens usque ad mare magnum et plaga ad meridiem australis 47:20 et plaga maris mare magnum a confi- 47:20 The west side also shall be the great sea from the

border, till a man come over against Hamath. This is the west side.

nio per directum donec uenias Emath haec est plaga maris 47:21 et diuidetis terram istam uobis per tribus Israhel 47:22 47:21 So shall ye divide this land unto you according to

the tribes of Israel.

47:22 And it shall come to pass, that ye shall divide it by lot for an inheritance unto you, and to the strangers that sojourn among you, which shall beget children among you: and they shall be unto you as born in the country among the children of Israel; they shall have inheritance with you among the tribes of Israel.

et mittetis eam in hereditatem uobis et aduenis qui accesserint ad uos qui genuerint filios in medio uestrum et erunt

Biblia Sacra Vulgata

uobis sicut indigenae inter filios Israhel uobiscum diuident

47:23 And it shall come to pass, that in what tribe the possessionem in medio tribuum Israhel 47:23 in tribu au-

stranger sojourneth, there shall ye give him his inheritance, saith the Lord GOD.

tem quacumque fuerit aduena ibi dabitis possessionem illi ait Dominus Deus

48:1 Now these are the names of the tribes. From the north 48:1 et haec nomina tribuum a finibus aquilonis iuxta uiam

end to the coast of the way of Hethlon, as one goeth to Hamath, Hazarenan, the border of Damascus northward, to the coast of Hamath; for these are his sides east and west; a portion for Dan.

Aethlon pergentibus Emath atrium Aenon terminus Damasci ad aquilonem iuxta Emath et erit ei plaga orientalis mare

48:2 And by the border of Dan, from the east side unto the Dan una 48:2 et ad terminum Dan a plaga orientali usque west side, a portion for Asher. 48:3 And by the border of Asher, from the east side even ad plagam maris Aser una 48:3 et super terminum Aser a unto the west side, a portion for Naphtali. 48:4 And by the border of Naphtali, from the east side unto plaga orientali usque ad plagam maris Nepthalim una 48:4 the west side, a portion for Manasseh.

et super terminum Nepthalim a plaga orientali usque ad pla-

48:5 And by the border of Manasseh, from the east side gam maris Manasse una 48:5 et super terminum Manasse a

unto the west side, a portion for Ephraim.

48:6 And by the border of Ephraim, from the east side plaga orientali usque ad plagam maris Ephraim una 48:6 et

even unto the west side, a portion for Reuben.

super terminum Ephraim a plaga orientali usque ad plagam

48:7 And by the border of Reuben, from the east side unto maris Ruben una 48:7 et super terminum Ruben a plaga the west side, a portion for Judah.

48:8 And by the border of Judah, from the east side unto orientali usque ad plagam maris Iuda una 48:8 et super ter-

the west side, shall be the offering which ye shall offer of five and twenty thousand reeds in breadth, and in length as one of the other parts, from the east side unto the west side: and the sanctuary shall be in the midst of it.

minum Iuda a plaga orientali usque ad plagam maris erunt primitiae quas separabitis uiginti quinque milibus latitudinis et longitudinis sicuti singulae partes a plaga orientali us-

48:9 The oblation that ye shall offer unto the LORD shall que ad plagam maris et erit sanctuarium in medio eius 48:9

be of five and twenty thousand in length, and of ten thousand in breadth.

primitiae quas separastis Domino longitudo uiginti quinque

48:10 And for them, even for the priests, shall be this holy milibus et latitudo decem milibus 48:10 hae autem erunt oblation; toward the north five and twenty thousand in length, and toward the west ten thousand in breadth, and toward the east ten thousand in breadth, and toward the south five and twenty thousand in length: and the sanctuary of the LORD shall be in the midst thereof.

primitiae sanctuarii sacerdotum ad aquilonem uiginti quinque milia et ad mare latitudinis decem milia sed et ad orientem latitudinis decem milia et ad meridiem longitudinis uiginti quinque milia et erit sanctuarium Domini in medio eius

48:11 It shall be for the priests that are sanctified of the 48:11 sacerdotibus sanctuarium erit de filiis Sadoc qui custo-

sons of Zadok; which have kept my charge, which went not astray when the children of Israel went astray, as the

Levites went astray. dierunt caerimonias meas et non errauerunt cum errarent filii 48:12 And this oblation of the land that is offered shall be Israhel sicut errauerunt et Leuitae 48:12 et erunt eis primi-

unto them a thing most holy by the border of the Levites.

tiae de primitiis terrae sanctum sanctorum iuxta terminum

48:13 And over against the border of the priests the Levites Leuitarum 48:13 sed et Leuitis similiter iuxta fines sacershall have five and twenty thousand in length, and ten thousand in breadth: all the length shall be five and twenty

thousand, and the breadth ten thousand. dotum uiginti quinque milia longitudinis et latitudinis decem milia omnis longitudo uiginti et quinque milium et latitudo

Prophetia Ezechielis

decem milium 48:14 et non uenundabunt ex eo neque muta- 48:14 And they shall not sell of it, neither exchange, nor alienate the firstfruits of the land: for it is holy unto the LORD.

bunt nec transferentur primitiae terrae quia sanctificatae sunt Domino 48:15 quinque milia autem quae supersunt in lati- 48:15 And the five thousand, that are left in the breadth

over against the five and twenty thousand, shall be a profane place for the city, for dwelling, and for suburbs: and

tudine per uiginti quinque milia profana erunt urbis in habi- the city shall be in the midst thereof. taculum et in suburbana et erit ciuitas in medio eius 48:16 48:16 And these shall be the measures thereof; the north

side four thousand and five hundred, and the south side four thousand and five hundred, and on the east side four thousand and five hundred, and the west side four thousand and five hundred.

et heae mensurae eius ad plagam septentrionalem quingenti et quattuor milia et ad plagam meridianam quingenti et quattuor milia et ad plagam orientalem quingenti et quattuor milia et ad plagam occidentalem quingenti et quattuor milia 48:17 erunt autem suburbana ciuitatis ad aquilonem ducenti 48:17 And the suburbs of the city shall be toward the north

two hundred and fifty, and toward the south two hundred and fifty, and toward the east two hundred and fifty, and

quinquaginta et in meridie ducenti quinquaginta et ad orien- toward the west two hundred and fifty. tem ducenti quinquaginta et ad mare ducenti quinquaginta 48:18 quod autem reliquum fuerit in longitudine secundum 48:18 And the residue in length over against the oblation

of the holy portion shall be ten thousand eastward, and ten thousand westward: and it shall be over against the oblation of the holy portion; and the increase thereof shall be for food unto them that serve the city.

primitias sanctuarii decem milia in orientem et decem milia ad occidentem erunt sicut primitiae sanctuarii et erunt fruges eius in panes his qui seruiunt ciuitati 48:19 seruientes au- 48:19 And they that serve the city shall serve it out of all

the tribes of Israel.

tem ciuitati operabuntur ex omnibus tribubus Israhel 48:20 48:20 All the oblation shall be five and twenty thousand by

five and twenty thousand: ye shall offer the holy oblation foursquare, with the possession of the city.

omnes primitiae uiginti quinque milium per uiginti quinque milia in quadrum separabuntur in primitias sanctuarii et possessionem ciuitatis 48:21 quod autem reliquum fuerit 48:21 And the residue shall be for the prince, on the one

side and on the other of the holy oblation, and of the possession of the city, over against the five and twenty thousand of the oblation toward the east border, and westward over against the five and twenty thousand toward the west border, over against the portions for the prince: and it shall be the holy oblation; and the sanctuary of the house shall be in the midst thereof.

principis erit ex omni parte primitiarum sanctuarii et possessionis ciuitatis e regione uiginti quinque milium primitiarum usque ad terminum orientalem sed et ad mare e regione uiginti quinque milium usque ad terminum maris similiter in partibus principis erit et erunt primitiae sanctuarii et sanctuarium templi in medio eius 48:22 de possessione au- 48:22 Moreover from the possession of the Levites, and

from the possession of the city, being in the midst of that which is the prince’s, between the border of Judah and the

tem Leuitarum et de possessione ciuitatis in medio partium border of Benjamin, shall be for the prince. principis erit inter terminum Iuda et inter terminum Beniamin et ad principem pertinebit 48:23 et reliquis tribubus a 48:23 As for the rest of the tribes, from the east side unto

the west side, Benjamin shall have a portion.

plaga orientali usque ad plagam occidentalem Beniamin una 48:24 et contra terminum Beniamin a plaga orientali usque ad 48:24 And by the border of Benjamin, from the east side

unto the west side, Simeon shall have a portion.

plagam occidentalem Symeon una 48:25 et super terminum 48:25 And by the border of Simeon, from the east side

unto the west side, Issachar a portion.

Symeonis a plaga orientali usque ad plagam occidentis Isa-

Biblia Sacra Vulgata

48:26 And by the border of Issachar, from the east side char una 48:26 et super terminum Isachar a plaga orientali

unto the west side, Zebulun a portion.

48:27 And by the border of Zebulun, from the east side usque ad plagam occidentalem Zabulon una 48:27 et super

unto the west side, Gad a portion.

terminum Zabulon a plaga orientali usque ad plagam maris

48:28 And by the border of Gad, at the south side south- Gad una 48:28 et super terminum Gad ad plagam austri in ward, the border shall be even from Tamar unto the waters of strife in Kadesh, and to the river toward the great sea.

meridiem et erit finis de Thamar usque ad aquas Contradic-

48:29 This is the land which ye shall divide by lot unto the tionis Cades hereditas contra mare magnum 48:29 haec est tribes of Israel for inheritance, and these are their portions, saith the Lord GOD.

terra quam mittetis in sortem tribubus Israhel et hae partitio-

48:30 And these are the goings out of the city on the north nes earum ait Dominus Deus 48:30 et hii egressus ciuitatis

side, four thousand and five hundred measures.

a plaga septentrionali quingentos et quattuor milia mensu-

48:31 And the gates of the city shall be after the names rabis 48:31 et portae ciuitatis in nominibus tribuum Isra-

of the tribes of Israel: three gates northward; one gate of Reuben, one gate of Judah, one gate of Levi.

hel portae tres a septentrione porta Ruben una porta Iudae

48:32 And at the east side four thousand and five hundred: una porta Leui una 48:32 et ad plagam orientalem quingen-

and three gates; and one gate of Joseph, one gate of Benjamin, one gate of Dan.

tos et quattuor milia et portae tres porta Ioseph una porta

48:33 And at the south side four thousand and five hundred Beniamin una porta Dan una 48:33 et ad plagam meridiameasures: and three gates; one gate of Simeon, one gate of Issachar, one gate of Zebulun.

nam quingentos et quattuor milia metieris portam Symeonis

48:34 At the west side four thousand and five hundred, unam portam Isachar unam portam Zabulon unam 48:34 et

with their three gates; one gate of Gad, one gate of Asher, one gate of Naphtali.

ad plagam occidentalem quingenti et quattuor milia portae eorum tres porta Gad una porta Aser una porta Nepthalim

48:35 It was round about eighteen thousand measures: and una 48:35 per circuitum decem et octo milia et nomen ciuithe name of the city from that day shall be, The LORD is there.

tatis ex illa die Dominus ibidem

Prophetia Danielis

1:1 anno tertio regni Ioachim regis Iuda uenit Nabuchodo- 1:1 In the third year of the reign of Jehoiakim king

of Judah came Nebuchadnezzar king of Babylon unto Jerusalem, and besieged it.

nosor rex Babylonis Hierusalem et obsedit eam 1:2 et tra-

1:2 And the Lord gave Jehoiakim king of Judah into his hand, with part of the vessels of the house of God: which he carried into the land of Shinar to the house of his god; and he brought the vessels into the treasure house of his god.

didit Dominus in manu eius Ioachim regem Iudae et partem uasorum domus Dei et asportauit ea in terram Sennaar in domum dei sui et uasa intulit in domum thesauri dei sui 1:3 et 1:3 And the king spake unto Ashpenaz the master of his

eunuchs, that he should bring certain of the children of Israel, and of the king’s seed, and of the princes;

ait rex Asfanaz praeposito eunuchorum suorum ut introduceret de filiis Israhel et de semine regio et tyrannorum 1:4 1:4 Children in whom was no blemish, but well favoured,

and skilful in all wisdom, and cunning in knowledge, and understanding science, and such as had ability in them to stand in the king’s palace, and whom they might teach the learning and the tongue of the Chaldeans.

pueros in quibus nulla esset macula decoros forma et eruditos omni sapientia cautos scientia et doctos disciplina et qui possent stare in palatio regis ut doceret eos litteras et linguam Chaldeorum 1:5 et constituit eis rex annonam per 1:5 And the king appointed them a daily provision of the

king’s meat, and of the wine which he drank: so nourishing them three years, that at the end thereof they might

singulos dies de cibis suis et de uino unde bibebat ipse ut stand before the king. enutriti tribus annis postea starent in conspectu regis 1:6 1:6 Now among these were of the children of Judah,

Daniel, Hananiah, Mishael, and Azariah:

fuerunt ergo inter eos de filiis Iuda Danihel Ananias Misahel et Azarias 1:7 et inposuit eis praepositus eunuchorum 1:7 Unto whom the prince of the eunuchs gave names:

for he gave unto Daniel the name of Belteshazzar; and to Hananiah, of Shadrach; and to Mishael, of Meshach; and

nomina Daniheli Balthasar et Ananiae Sedrac Misaheli Mi- to Azariah, of Abednego. sac et Azariae Abdenago 1:8 proposuit autem Danihel in 1:8 But Daniel purposed in his heart that he would not

defile himself with the portion of the king’s meat, nor with the wine which he drank: therefore he requested of the

corde suo ne pollueretur de mensa regis neque de uino po- prince of the eunuchs that he might not defile himself. tus eius et rogauit eunuchorum praepositum ne contaminaretur 1:9 dedit autem Deus Daniheli gratiam et misericor- 1:9 Now God had brought Daniel into favour and tender

love with the prince of the eunuchs.

diam in conspectu principis eunuchorum 1:10 et ait prin- 1:10 And the prince of the eunuchs said unto Daniel, I fear

my lord the king, who hath appointed your meat and your drink: for why should he see your faces worse liking than the children which are of your sort? then shall ye make me endanger my head to the king.

ceps eunuchorum ad Danihel timeo ego dominum meum regem qui constituit uobis cibum et potum qui si uiderit uultus uestros macilentiores prae ceteris adulescentibus coaeuis uestris condemnabitis caput meum regi 1:11 et dixit Danihel 1:11 Then said Daniel to Melzar, whom the prince of

the eunuchs had set over Daniel, Hananiah, Mishael, and Azariah,

ad Malassar quem constituerat princeps eunuchorum super

Danihel Ananiam Misahel et Azariam 1:12 tempta nos ob- 1:12 Prove thy servants, I beseech thee, ten days; and let

them give us pulse to eat, and water to drink.

secro seruos tuos diebus decem et dentur nobis legumina ad uescendum et aqua ad bibendum 1:13 et contemplare uul- 1:13 Then let our countenances be looked upon before

thee, and the countenance of the children that eat of the portion of the king’s meat: and as thou seest, deal with thy

tus nostros et uultus puerorum qui uescuntur cibo regio et servants. sicut uideris facies cum seruis tuis 1:14 qui audito sermone 1:14 So he consented to them in this matter, and proved

them ten days.

huiuscemodi temptauit eos diebus decem 1:15 post dies au- 1:15 And at the end of ten days their countenances ap-

peared fairer and fatter in flesh than all the children which did eat the portion of the king’s meat.

tem decem apparuerunt uultus eorum meliores et corpulen-

Biblia Sacra Vulgata

1:16 Thus Melzar took away the portion of their meat, and tiores prae omnibus pueris qui uescebantur cibo regio 1:16

the wine that they should drink; and gave them pulse.

porro Malassar tollebat cibaria et uinum potus eorum dabat-

1:17 As for these four children, God gave them knowl- que eis legumina 1:17 pueris autem his dedit Deus scien-

edge and skill in all learning and wisdom: and Daniel had understanding in all visions and dreams.

tiam et disciplinam in omni libro et sapientia Daniheli au-

1:18 Now at the end of the days that the king had said tem intellegentiam omnium uisionum et somniorum 1:18

he should bring them in, then the prince of the eunuchs brought them in before Nebuchadnezzar.

conpletis itaque diebus post quos dixerat rex ut introducerentur introduxit eos praepositus eunuchorum in conspectu

1:19 And the king communed with them; and among them Nabuchodonosor 1:19 cumque locutus eis fuisset rex non

all was found none like Daniel, Hananiah, Mishael, and Azariah: therefore stood they before the king.

sunt inuenti de uniuersis tales ut Danihel Ananias Misahel

1:20 And in all matters of wisdom and understanding, that et Azarias et steterunt in conspectu regis 1:20 et omne uer-

the king enquired of them, he found them ten times better than all the magicians and astrologers that were in all his

realm. bum sapientiae et intellectus quod sciscitatus est ab eis rex

inuenit in eis decuplum super cunctos ariolos et magos qui

1:21 And Daniel continued even unto the first year of king erant in uniuerso regno eius 1:21 fuit autem Danihel usque

Cyrus.

ad annum primum Cyri regis

2:1 And in the second year of the reign of Nebuchadnez- 2:1 in anno secundo regni Nabuchodonosor uidit Nabuzar Nebuchadnezzar dreamed dreams, wherewith his spirit was troubled, and his sleep brake from him.

chodonosor somnium et conterritus est spiritus eius et som-

2:2 Then the king commanded to call the magicians, and nium eius fugit ab eo 2:2 praecepit ergo rex ut conuoca-

the astrologers, and the sorcerers, and the Chaldeans, for to shew the king his dreams. So they came and stood be-

fore the king. rentur arioli et magi et malefici et Chaldei et indicarent re-2:3 And the king said unto them, I have dreamed a dream, gi somnia sua qui cum uenissent steterunt coram rege 2:3

and my spirit was troubled to know the dream.

et dixit ad eos rex uidi somnium et mente confusus ignoro

2:4 Then spake the Chaldeans to the king in Syriack, O quid uiderim 2:4 responderuntque Chaldei regi syriace rex

king, live for ever: tell thy servants the dream, and we will shew the interpretation.

in sempiternum uiue dic somnium seruis tuis et interpreta-

2:5 The king answered and said to the Chaldeans, The tionem eius indicabimus 2:5 et respondens rex ait Chal-

thing is gone from me: if ye will not make known unto me the dream, with the interpretation thereof, ye shall be

cut in pieces, and your houses shall be made a dunghill. deis sermo recessit a me nisi indicaueritis mihi somnium et

coniecturam eius peribitis uos et domus uestrae publicabun-

2:6 But if ye shew the dream, and the interpretation tur 2:6 si autem somnium et coniecturam eius narraueritis

thereof, ye shall receive of me gifts and rewards and great honour: therefore shew me the dream, and the interpreta-

tion thereof. praemia et dona et honorem multum accipietis a me som-2:7 They answered again and said, Let the king tell his nium igitur et interpretationem eius indicate mihi 2:7 re-

servants the dream, and we will shew the interpretation of it.

sponderunt secundo atque dixerunt rex somnium dicat ser-

2:8 The king answered and said, I know of certainty that uis suis et interpretationem illius indicabimus 2:8 respon-

ye would gain the time, because ye see the thing is gone from me.

dit rex et ait certo noui quia tempus rediiitis scientes quod

2:9 But if ye will not make known unto me the dream, recesserit a me sermo 2:9 si ergo somnium non indicaue-

there is but one decree for you: for ye have prepared lying and corrupt words to speak before me, till the time be changed: therefore tell me the dream, and I shall know that ye can shew me the interpretation thereof.

ritis mihi una est de uobis sententia quod interpretationem

Prophetia Danielis

quoque fallacem et deceptione plenam conposueritis ut loquamini mihi donec tempus pertranseat somnium itaque dicite mihi ut sciam quod interpretationem quoque eius ueram loquamini 2:10 respondentes ergo Chaldei coram rege 2:10 The Chaldeans answered before the king, and said,

There is not a man upon the earth that can shew the king’s matter: therefore there is no king, lord, nor ruler, that asked such things at any magician, or astrologer, or Chaldean.

dixerunt non est homo super terram qui sermonem tuum rex possit implere sed neque regum quisquam magnus et potens uerbum huiuscemodi sciscitatur ab omni ariolo et mago et

Chaldeo 2:11 sermo enim quem tu rex quaeris grauis est 2:11 And it is a rare thing that the king requireth, and there is none other that can shew it before the king, except the gods, whose dwelling is not with flesh.

nec repperietur quisquam qui indicet illum in conspectu regis exceptis diis quorum non est cum hominibus conuersatio 2:12 quo audito rex in furore et in ira magna praecepit ut 2:12 For this cause the king was angry and very furious,

and commanded to destroy all the wise men of Babylon.

perirent omnes sapientes Babylonis 2:13 et egressa senten- 2:13 And the decree went forth that the wise men should

be slain; and they sought Daniel and his fellows to be slain.

tia sapientes interficiebantur quaerebaturque Danihel et socii eius ut perirent 2:14 tunc Danihel requisiuit de lege atque 2:14 Then Daniel answered with counsel and wisdom to

Arioch the captain of the king’s guard, which was gone forth to slay the wise men of Babylon:

sententia ab Arioch principe militiae regis qui egressus fuerat ad interficiendos sapientes Babylonis 2:15 et interroga- 2:15 He answered and said to Arioch the king’s captain,

Why is the decree so hasty from the king? Then Arioch made the thing known to Daniel.

uit eum qui a rege acceperat potestatem quam ob causam tam crudelis sententia a facie esset regis egressa cum ergo rem indicasset Arioch Daniheli 2:16 Danihel ingressus ro- 2:16 Then Daniel went in, and desired of the king that he

would give him time, and that he would shew the king the interpretation.

gauit regem ut tempus daret sibi ad solutionem indicandam regi 2:17 et ingressus est domum suam Ananiaeque Misa- 2:17 Then Daniel went to his house, and made the thing

known to Hananiah, Mishael, and Azariah, his companions:

heli et Azariae sociis suis indicauit negotium 2:18 ut quae-

2:18 That they would desire mercies of the God of heaven concerning this secret; that Daniel and his fellows should not perish with the rest of the wise men of Babylon.

rerent misericordiam a facie Dei caeli super sacramento isto et non perirent Danihel et socii eius cum ceteris sapientibus Babylonis 2:19 tunc Daniheli per uisionem nocte my- 2:19 Then was the secret revealed unto Daniel in a night

vision. Then Daniel blessed the God of heaven.

sterium reuelatum est et Danihel benedixit Deo caeli 2:20 2:20 Daniel answered and said, Blessed be the name of

God for ever and ever: for wisdom and might are his:

et locutus ait sit nomen Domini benedictum a saeculo et usque in saeculum quia sapientia et fortitudo eius sunt 2:21 et 2:21 And he changeth the times and the seasons: he re-

moveth kings, and setteth up kings: he giveth wisdom unto the wise, and knowledge to them that know understanding:

ipse mutat tempora et aetates transfert regna atque constituit dat sapientiam sapientibus et scientiam intellegentibus disciplinam 2:22 ipse reuelat profunda et abscondita et nouit in 2:22 He revealeth the deep and secret things: he knoweth

what is in the darkness, and the light dwelleth with him.

tenebris constituta et lux cum eo est 2:23 tibi Deus patrum 2:23 I thank thee, and praise thee, O thou God of my

fathers, who hast given me wisdom and might, and hast made known unto me now what we desired of thee: for

meorum confiteor teque laudo quia sapientiam et fortitudi- thou hast now made known unto us the king’s matter.

Biblia Sacra Vulgata

nem dedisti mihi et nunc ostendisti mihi quae rogauimus te

2:24 Therefore Daniel went in unto Arioch, whom the quia sermonem regis aperuisti nobis 2:24 post haec Dani-

king had ordained to destroy the wise men of Babylon: he went and said thus unto him; Destroy not the wise men of Babylon: bring me in before the king, and I will shew unto the king the interpretation.

hel ingressus ad Arioch quem constituerat rex ut perderet sapientes Babylonis sic ei locutus est sapientes Babylonis ne perdas introduc me in conspectu regis et solutionem re-

2:25 Then Arioch brought in Daniel before the king in gi enarrabo 2:25 tunc Arioch festinus introduxit Danihelem

haste, and said thus unto him, I have found a man of the captives of Judah, that will make known unto the king the

interpretation. ad regem et dixit ei inueni hominem de filiis transmigratio-2:26 The king answered and said to Daniel, whose name nis Iudae qui solutionem regi adnuntiet 2:26 respondit rex

was Belteshazzar, Art thou able to make known unto me the dream which I have seen, and the interpretation

thereof? et dixit Daniheli cuius nomen erat Balthasar putasne uere po-

tes indicare mihi somnium quod uidi et interpretationem eius

2:27 Daniel answered in the presence of the king, and said, 2:27 et respondens Danihel coram rege ait mysterium quod The secret which the king hath demanded cannot the wise men, the astrologers, the magicians, the soothsayers, shew

unto the king; rex interrogat sapientes magi et arioli et aruspices non que-2:28 But there is a God in heaven that revealeth secrets, unt indicare regi 2:28 sed est Deus in caelo reuelans myste-

and maketh known to the king Nebuchadnezzar what shall be in the latter days. Thy dream, and the visions of thy

head upon thy bed, are these; ria qui indicauit tibi rex Nabuchodonosor quae uentura sunt

nouissimis temporibus somnium tuum et uisiones capitis tui

2:29 As for thee, O king, thy thoughts came into thy mind in cubili tuo huiuscemodi sunt 2:29 tu rex cogitare coepisti

upon thy bed, what should come to pass hereafter: and he that revealeth secrets maketh known to thee what shall

come to pass. in stratu tuo quid esset futurum post haec et qui reuelat my-2:30 But as for me, this secret is not revealed to me for steria ostendit tibi quae uentura sunt 2:30 mihi quoque non

any wisdom that I have more than any living, but for their sakes that shall make known the interpretation to the king,

and that thou mightest know the thoughts of thy heart. in sapientia quae est in me plus quam in cunctis uiuentibus

sacramentum hoc reuelatum est sed ut interpretatio regi ma-

2:31 Thou, O king, sawest, and behold a great image. This nifesta fieret et cogitationes mentis tuae scires 2:31 tu rex

great image, whose brightness was excellent, stood before thee; and the form thereof was terrible.

uidebas et ecce quasi statua una grandis statua illa magna et statura sublimis stabat contra te et intuitus eius erat terribilis

2:32 This image’s head was of fine gold, his breast and his 2:32 huius statuae caput ex auro optimo erat pectus autem et

arms of silver, his belly and his thighs of brass, 2:33 His legs of iron, his feet part of iron and part of clay. brachia de argento porro uenter et femora ex aere 2:33 ti-

biae autem ferreae pedum quaedam pars erat ferrea quaedam

2:34 Thou sawest till that a stone was cut out without fictilis 2:34 uidebas ita donec abscisus est lapis sine mani-

hands, which smote the image upon his feet that were of iron and clay, and brake them to pieces.

bus et percussit statuam in pedibus eius ferreis et fictilibus et

2:35 Then was the iron, the clay, the brass, the silver, and comminuit eos 2:35 tunc contrita sunt pariter ferrum testa

the gold, broken to pieces together, and became like the chaff of the summer threshingfloors; and the wind carried them away, that no place was found for them: and the stone that smote the image became a great mountain, and filled the whole earth.

aes argentum et aurum et redacta quasi in fauillam aestiuae areae rapta sunt uento nullusque locus inuentus est eis lapis autem qui percusserat statuam factus est mons magnus et im-

2:36 This is the dream; and we will tell the interpretation pleuit uniuersam terram 2:36 hoc est somnium interpretathereof before the king.

Prophetia Danielis

tionem quoque eius dicemus coram te rex 2:37 tu rex regum 2:37 Thou, O king, art a king of kings: for the God of

heaven hath given thee a kingdom, power, and strength, and glory.

es et Deus caeli regnum fortitudinem et imperium et gloriam dedit tibi 2:38 et omnia in quibus habitant filii hominum et 2:38 And wheresoever the children of men dwell, the

beasts of the field and the fowls of the heaven hath he given into thine hand, and hath made thee ruler over them

bestiae agri uolucresque caeli dedit in manu tua et sub di- all. Thou art this head of gold.

cione tua uniuersa constituit tu es ergo caput aureum 2:39 2:39 And after thee shall arise another kingdom inferior to thee, and another third kingdom of brass, which shall bear rule over all the earth.

et post te consurget regnum aliud minus te et regnum tertium aliud aereum quod imperabit uniuersae terrae 2:40 et reg- 2:40 And the fourth kingdom shall be strong as iron: foras-

much as iron breaketh in pieces and subdueth all things: and as iron that breaketh all these, shall it break in pieces

num quartum erit uelut ferrum quomodo ferrum comminuit and bruise. et domat omnia sic comminuet omnia haec et conteret 2:41 2:41 And whereas thou sawest the feet and toes, part of

potters’ clay, and part of iron, the kingdom shall be divided; but there shall be in it of the strength of the iron,

porro quia uidisti pedum et digitorum partem testae figuli et forasmuch as thou sawest the iron mixed with miry clay. partem ferream regnum diuisum erit quod tamen de plantario ferri orietur secundum quod uidisti ferrum mixtum testae ex luto 2:42 et digitos pedum ex parte ferreos et ex parte 2:42 And as the toes of the feet were part of iron, and part

of clay, so the kingdom shall be partly strong, and partly broken.

fictiles ex parte regnum erit solidum et ex parte contritum 2:43 quia autem uidisti ferrum mixtum testae ex luto commi- 2:43 And whereas thou sawest iron mixed with miry clay,

they shall mingle themselves with the seed of men: but they shall not cleave one to another, even as iron is not

scebuntur quidem humano semine sed non adherebunt sibi mixed with clay. sicuti ferrum misceri non potest testae 2:44 in diebus au- 2:44 And in the days of these kings shall the God of

heaven set up a kingdom, which shall never be destroyed: and the kingdom shall not be left to other people, but it shall break in pieces and consume all these kingdoms, and it shall stand for ever.

tem regnorum illorum suscitabit Deus caeli regnum quod in aeternum non dissipabitur et regnum eius populo alteri non tradetur comminuet et consumet uniuersa regna haec et ipsum stabit in aeternum 2:45 secundum quod uidisti quod de 2:45 Forasmuch as thou sawest that the stone was cut out

of the mountain without hands, and that it brake in pieces the iron, the brass, the clay, the silver, and the gold; the great God hath made known to the king what shall come to pass hereafter: and the dream is certain, and the interpretation thereof sure.

monte abscisus est lapis sine manibus et comminuit testam et ferrum et aes et argentum et aurum Deus magnus ostendit regi quae futura sunt postea et uerum est somnium et fidelis interpretatio eius 2:46 tunc rex Nabuchodonosor cecidit in 2:46 Then the king Nebuchadnezzar fell upon his face, and

worshipped Daniel, and commanded that they should offer an oblation and sweet odours unto him.

faciem suam et Danihelum adorauit et hostias et incensum praecepit ut sacrificarent ei 2:47 loquens ergo rex ait Dani- 2:47 The king answered unto Daniel, and said, Of a truth

it is, that your God is a God of gods, and a Lord of kings, and a revealer of secrets, seeing thou couldest reveal this

heli uere Deus uester Deus deorum est et Dominus regum secret. et reuelans mysteria quoniam potuisti aperire sacramentum hoc 2:48 tunc rex Danihelum in sublime extulit et munera 2:48 Then the king made Daniel a great man, and gave

him many great gifts, and made him ruler over the whole province of Babylon, and chief of the governors over all

multa et magna dedit ei et constituit eum principem super the wise men of Babylon. omnes prouincias Babylonis et praefectum magistratuum super cunctos sapientes Babylonis 2:49 Danihel autem post- 2:49 Then Daniel requested of the king, and he set

Shadrach, Meshach, and Abednego, over the affairs of the province of Babylon: but Daniel sat in the gate of the king.

Biblia Sacra Vulgata

ulauit a rege et constituit super opera prouinciae Babylonis Sedrac Misac et Abdenago ipse autem Danihel erat in foribus regis

3:1 Nebuchadnezzar the king made an image of gold, 3:1 Nabuchodonosor rex fecit statuam auream altitudine

whose height was threescore cubits, and the breadth thereof six cubits: he set it up in the plain of Dura, in the

province of Babylon. cubitorum sexaginta latitudine cubitorum sex et statuit eam

3:2 Then Nebuchadnezzar the king sent to gather together in campo Duram prouinciae Babylonis 3:2 itaque Nabuthe princes, the governors, and the captains, the judges, the treasurers, the counsellors, the sheriffs, and all the rulers of the provinces, to come to the dedication of the image which Nebuchadnezzar the king had set up.

chodonosor rex misit ad congregandos satrapas magistratus et iudices duces et tyrannos et praefectos omnesque principes regionum ut conuenirent ad dedicationem statuae quam

3:3 Then the princes, the governors, and captains, the erexerat Nabuchodonosor rex 3:3 tunc congregati sunt sa-

judges, the treasurers, the counsellors, the sheriffs, and all the rulers of the provinces, were gathered together unto the dedication of the image that Nebuchadnezzar the king had set up; and they stood before the image that Nebuchadnezzar had set up.

trapae magistratus et iudices duces et tyranni et optimates qui erant in potestatibus constituti et uniuersi principes regionum ut conuenirent ad dedicationem statuae quam erexerat Nabuchodonosor rex stabant autem in conspectu statuae

3:4 Then an herald cried aloud, To you it is commanded, quam posuerat Nabuchodonosor 3:4 et praeco clamabat ua-

O people, nations, and languages,

3:5 That at what time ye hear the sound of the cornet, flute, lenter uobis dicitur populis tribubus et linguis 3:5 in hora

harp, sackbut, psaltery, dulcimer, and all kinds of musick, ye fall down and worship the golden image that Nebuchad-

nezzar the king hath set up: qua audieritis sonitum tubae et fistulae et citharae sambucae et psalterii et symphoniae et uniuersi generis musicorum cadentes adorate statuam auream quam constituit Nabuchodo-

3:6 And whoso falleth not down and worshippeth shall the nosor rex 3:6 si quis autem non prostratus adorauerit eadem

same hour be cast into the midst of a burning fiery furnace.

3:7 Therefore at that time, when all the people heard the hora mittetur in fornacem ignis ardentis 3:7 post haec igi-

sound of the cornet, flute, harp, sackbut, psaltery, and all kinds of musick, all the people, the nations, and the languages, fell down and worshipped the golden image that Nebuchadnezzar the king had set up.

tur statim ut audierunt omnes populi sonitum tubae fistulae et citharae sambucae et psalterii et symphoniae et omnis generis musicorum cadentes omnes populi et tribus et linguae adorauerunt statuam auream quam constituerat Nabuchodo-

3:8 Wherefore at that time certain Chaldeans came near, nosor rex 3:8 statimque et in ipso tempore accedentes uiri

and accused the Jews.

3:9 They spake and said to the king Nebuchadnezzar, O chaldei accusauerunt Iudaeos 3:9 dixeruntque Nabuchodo-

king, live for ever.

3:10 Thou, O king, hast made a decree, that every man nosor regi rex in aeternum uiue 3:10 tu rex posuisti decre-

that shall hear the sound of the cornet, flute, harp, sackbut, psaltery, and dulcimer, and all kinds of musick, shall fall

down and worship the golden image: tum ut omnis homo qui audierit sonitum tubae fistulae et ci-

tharae sambucae et psalterii et symphoniae et uniuersi gene-

3:11 And whoso falleth not down and worshippeth, that he ris musicorum prosternat se et adoret statuam auream 3:11

should be cast into the midst of a burning fiery furnace.

si quis autem non procidens adorauerit mittatur in fornacem

3:12 There are certain Jews whom thou hast set over the ignis ardentem 3:12 sunt ergo uiri iudaei quos constituisti

affairs of the province of Babylon, Shadrach, Meshach, and Abednego; these men, O king, have not regarded thee: they serve not thy gods, nor worship the golden image which thou hast set up.

Prophetia Danielis

super opera regionis Babyloniae Sedrac Misac et Abdenago uiri isti contempserunt rex decretum tuum deos tuos non colunt et statuam auream quam erexisti non adorant 3:13 tunc 3:13 Then Nebuchadnezzar in his rage and fury com-

manded to bring Shadrach, Meshach, and Abednego. Then they brought these men before the king.

Nabuchodonosor in furore et in ira praecepit ut adducerentur Sedrac Misac et Abdenago qui confestim adducti sunt in conspectu regis 3:14 pronuntiansque Nabuchodonosor rex 3:14 Nebuchadnezzar spake and said unto them, Is it true,

O Shadrach, Meshach, and Abednego, do not ye serve my gods, nor worship the golden image which I have set up?

ait eis uerene Sedrac Misac et Abdenago deos meos non colitis et statuam auream quam constitui non adoratis 3:15 3:15 Now if ye be ready that at what time ye hear the sound

of the cornet, flute, harp, sackbut, psaltery, and dulcimer, and all kinds of musick, ye fall down and worship the image which I have made; well: but if ye worship not, ye shall be cast the same hour into the midst of a burning fiery furnace; and who is that God that shall deliver you out of my hands?

nunc ergo si estis parati quacumque hora audieritis sonitum tubae fistulae et citharae sambucae psalterii et symphoniae omnisque generis musicorum prosternite uos et adorate statuam quam feci quod si non adoraueritis eadem hora mittemini in fornacem ignis ardentem et quis est Deus qui eripiat uos de manu mea 3:16 respondentes Sedrac Misac et Ab- 3:16 Shadrach, Meshach, and Abednego, answered and

said to the king, O Nebuchadnezzar, we are not careful to answer thee in this matter.

denago dixerunt regi Nabuchodonosor non oportet nos de hac re respondere tibi 3:17 ecce enim Deus noster quem 3:17 If it be so, our God whom we serve is able to deliver

us from the burning fiery furnace, and he will deliver us out of thine hand, O king.

colimus potest eripere nos de camino ignis ardentis et de manibus tuis rex liberare 3:18 quod si noluerit notum tibi 3:18 But if not, be it known unto thee, O king, that we will

not serve thy gods, nor worship the golden image which thou hast set up.

sit rex quia deos tuos non colimus et statuam auream quam erexisti non adoramus 3:19 tunc Nabuchodonosor repletus 3:19 Then was Nebuchadnezzar full of fury, and the form

of his visage was changed against Shadrach, Meshach, and Abednego: therefore he spake, and commanded that they should heat the furnace one seven times more than it was wont to be heated.

est furore et aspectus faciei illius inmutatus est super Sedrac Misac et Abdenago et praecepit ut succenderetur fornax septuplum quam succendi consuerat 3:20 et uiris fortissi- 3:20 And he commanded the most mighty men that were

in his army to bind Shadrach, Meshach, and Abednego, and to cast them into the burning fiery furnace.

mis de exercitu suo iussit ut ligatis pedibus Sedrac Misac et Abdenago mitterent eos in fornacem ignis ardentem 3:21 3:21 Then these men were bound in their coats, their ho-

sen, and their hats, and their other garments, and were cast into the midst of the burning fiery furnace.

et confestim uiri illi uincti cum bracis suis et tiaris et calciamentis et uestibus missi sunt in medium fornacis ignis ardentis 3:22 nam iussio regis urguebat fornax autem suc- 3:22 Therefore because the king’s commandment was ur-

gent, and the furnace exceeding hot, the flames of the fire slew those men that took up Shadrach, Meshach, and

censa erat nimis porro uiros illos qui miserant Sedrac Misac Abednego. et Abdenago interfecit flamma ignis 3:23 uiri autem hii id 3:23 And these three men, Shadrach, Meshach, and Abed-

nego, fell down bound into the midst of the burning fiery furnace.

est tres Sedrac Misac et Abdenago ceciderunt in medio camini ignis ardentis conligati

Biblia Sacra Vulgata

MCCLX

Prophetia Osee

1:1 uerbum Domini quod factum est ad Osee filium Be- 1:1 The word of the LORD that came unto Hosea, the

son of Beeri, in the days of Uzziah, Jotham, Ahaz, and Hezekiah, kings of Judah, and in the days of Jeroboam the

eri in diebus Oziae Ioatham Ahaz Ezechiae regum Iuda et son of Joash, king of Israel. in diebus Hieroboam filii Ioas regis Israhel 1:2 principium 1:2 The beginning of the word of the LORD by Hosea.

And the LORD said to Hosea, Go, take unto thee a wife of whoredoms and children of whoredoms: for the land hath

loquendi Dominum in Osee et dixit Dominus ad Osee uade committed great whoredom, departing from the LORD. sume tibi uxorem fornicationum et filios fornicationum quia fornicans fornicabitur terra a Domino 1:3 et abiit et accepit 1:3 So he went and took Gomer the daughter of Diblaim;

which conceived, and bare him a son.

Gomer filiam Debelaim et concepit et peperit filium 1:4 et 1:4 And the LORD said unto him, Call his name Jezreel;

for yet a little while, and I will avenge the blood of Jezreel upon the house of Jehu, and will cause to cease the king-

dixit Dominus ad eum uoca nomen eius Hiezrahel quoniam dom of the house of Israel. adhuc modicum et uisitabo sanguinem Hiezrahel super domum Hieu et quiescere faciam regnum domus Israhel 1:5 et 1:5 And it shall come to pass at that day, that I will break

the bow of Israel, in the valley of Jezreel.

in illa die conteram arcum Israhel in ualle Hiezrahel 1:6 et 1:6 And she conceived again, and bare a daughter. And

God said unto him, Call her name Loruhamah: for I will no more have mercy upon the house of Israel; but I will

concepit adhuc et peperit filiam et dixit ei uoca nomen eius utterly take them away. Absque misericordia quia non addam ultra misereri domui Israhel sed obliuione obliuiscar eorum 1:7 et domui Iuda 1:7 But I will have mercy upon the house of Judah, and

will save them by the LORD their God, and will not save them by bow, nor by sword, nor by battle, by horses, nor

miserebor et saluabo eos in Domino Deo suo et non saluabo by horsemen. eos in arcu et gladio et in bello et in equis et in equitibus 1:8 1:8 Now when she had weaned Loruhamah, she con-

ceived, and bare a son.

et ablactauit eam quae erat absque misericordia et concepit et peperit filium 1:9 et dixit uoca nomen eius Non populus 1:9 Then said God, Call his name Loammi: for ye are not

my people, and I will not be your God.

meus quia uos non populus meus et ego non ero uester 1:10 1:10 Yet the number of the children of Israel shall be as the sand of the sea, which cannot be measured nor numbered; and it shall come to pass, that in the place where it was said unto them, Ye are not my people, there it shall be said unto them, Ye are the sons of the living God.

et erit numerus filiorum Israhel quasi harena maris quae sine mensura est et non numerabitur et erit in loco ubi dicetur eis non populus meus uos dicetur eis filii Dei uiuentis 1:11 1:11 Then shall the children of Judah and the children of

Israel be gathered together, and appoint themselves one head, and they shall come up out of the land: for great

et congregabuntur filii Iuda et filii Israhel pariter et ponent shall be the day of Jezreel. sibimet caput unum et ascendent de terra quia magnus dies Hiezrahel

2:1 dicite fratribus uestris Populus meus et sorori uestrae 2:1 Say ye unto your brethren, Ammi; and to your sisters,

Ruhamah.

Misericordiam consecuta 2:2 iudicate matrem uestram iu- 2:2 Plead with your mother, plead: for she is not my wife,

neither am I her husband: let her therefore put away her whoredoms out of her sight, and her adulteries from be-

dicate quoniam ipsa non uxor mea et ego non uir eius auferat tween her breasts; fornicationes suas a facie sua et adulteria sua de medio ube-

rum suorum 2:3 ne forte expoliem eam nudam et statuam 2:3 Lest I strip her naked, and set her as in the day that she was born, and make her as a wilderness, and set her like a dry land, and slay her with thirst.

eam secundum diem natiuitatis suae et ponam eam quasi solitudinem et statuam eam uelut terram inuiam et interficiam eam siti 2:4 et filiorum illius non miserebor quoniam fi- 2:4 And I will not have mercy upon her children; for they

be the children of whoredoms.

Biblia Sacra Vulgata

2:5 For their mother hath played the harlot: she that con- lii fornicationum sunt 2:5 quia fornicata est mater eorum

ceived them hath done shamefully: for she said, I will go after my lovers, that give me my bread and my water, my

wool and my flax, mine oil and my drink. confusa est quae concepit eos quia dixit uadam post ama-

tores meos qui dant panes mihi et aquas meas lanam meam

2:6 Therefore, behold, I will hedge up thy way with thorns, et linum meum oleum meum et potum meum 2:6 propter and make a wall, that she shall not find her paths.

hoc ecce ego sepiam uiam tuam spinis et sepiam eam ma-

2:7 And she shall follow after her lovers, but she shall not ceria et semitas suas non inueniet 2:7 et sequetur amatoovertake them; and she shall seek them, but shall not find them: then shall she say, I will go and return to my first

husband; for then was it better with me than now. res suos et non adprehendet eos et quaeret eos et non inue-

niet et dicet uadam et reuertar ad uirum meum priorem quia

2:8 For she did not know that I gave her corn, and wine, bene mihi erat tunc magis quam nunc 2:8 et haec nesciand oil, and multiplied her silver and gold, which they prepared for Baal.

uit quia ego dedi ei frumentum et uinum et oleum et argen-

2:9 Therefore will I return, and take away my corn in the tum multiplicaui ei et aurum quae fecerunt Baal 2:9 idcirco

time thereof, and my wine in the season thereof, and will recover my wool and my flax given to cover her naked-

ness. conuertar et sumam frumentum meum in tempore suo et ui-

num meum in tempore suo et liberabo lanam meam et linum

2:10 And now will I discover her lewdness in the sight of meum quae operiebant ignominiam eius 2:10 et nunc reue-

her lovers, and none shall deliver her out of mine hand.

labo stultitiam eius in oculis amatorum eius et uir non eruet

2:11 I will also cause all her mirth to cease, her feast eam de manu mea 2:11 et cessare faciam omne gaudium days, her new moons, and her sabbaths, and all her solemn feasts.

eius sollemnitatem eius neomeniam eius sabbatum eius et

2:12 And I will destroy her vines and her fig trees, whereof omnia festa tempora eius 2:12 et corrumpam uineam eius she hath said, These are my rewards that my lovers have given me: and I will make them a forest, and the beasts of

the field shall eat them. et ficum eius de quibus dixit mercedes hae meae sunt quas

dederunt mihi amatores mei et ponam eam in saltu et come-

2:13 And I will visit upon her the days of Baalim, wherein det illam bestia agri 2:13 et uisitabo super eam dies Baalim

she burned incense to them, and she decked herself with her earrings and her jewels, and she went after her lovers,

and forgat me, saith the LORD. quibus accendebat incensum et ornabatur inaure sua et mo-

nili suo et ibat post amatores suos et mei obliuiscebatur dicit

2:14 Therefore, behold, I will allure her, and bring her into Dominus 2:14 propter hoc ecce ego lactabo eam et ducam the wilderness, and speak comfortably unto her.

2:15 And I will give her her vineyards from thence, and eam in solitudinem et loquar ad cor eius 2:15 et dabo ei ui-

the valley of Achor for a door of hope: and she shall sing there, as in the days of her youth, and as in the day when

she came up out of the land of Egypt. nitores eius ex eodem loco et uallem Achor ad aperiendam

spem et canet ibi iuxta dies iuuentutis suae et iuxta dies as-

2:16 And it shall be at that day, saith the LORD, that thou censionis suae de terra Aegypti 2:16 et erit in die illo ait

shalt call me Ishi; and shalt call me no more Baali.

Dominus uocabit me Vir meus et non uocabit me ultra Baali

2:17 For I will take away the names of Baalim out of her 2:17 et auferam nomina Baalim de ore eius et non recordabi-

mouth, and they shall no more be remembered by their name.

tur ultra nominis eorum 2:18 et percutiam eis foedus in die

2:18 And in that day will I make a covenant for them with the beasts of the field and with the fowls of heaven, and with the creeping things of the ground: and I will break the bow and the sword and the battle out of the earth, and will make them to lie down safely.

illa cum bestia agri et cum uolucre caeli et cum reptili terrae et arcum et gladium et bellum conteram de terra et dormire

Prophetia Osee

eos faciam fiducialiter 2:19 et sponsabo te mihi in sempi- 2:19 And I will betroth thee unto me for ever; yea, I will

betroth thee unto me in righteousness, and in judgment, and in lovingkindness, and in mercies.

ternum et sponsabo te mihi in iustitia et iudicio et in misericordia et miserationibus 2:20 et sponsabo te mihi in fide et 2:20 I will even betroth thee unto me in faithfulness: and

thou shalt know the LORD.

scies quia ego Dominus 2:21 et erit in illa die exaudiam di- 2:21 And it shall come to pass in that day, I will hear, saith

the LORD, I will hear the heavens, and they shall hear the earth;

cit Dominus exaudiam caelos et illi exaudient terram 2:22

2:22 And the earth shall hear the corn, and the wine, and the oil; and they shall hear Jezreel.

et terra exaudiet triticum et uinum et oleum et haec exaudient Hiezrahel 2:23 et seminabo eam mihi in terram et misere- 2:23 And I will sow her unto me in the earth; and I will

have mercy upon her that had not obtained mercy;

bor eius quae fuit absque misericordia 2:24 et dicam non 2:24 and I will say to them which were not my people,

Thou art my people; and they shall say, Thou art my God.

populo meo populus meus tu et ipse dicet Dominus meus es tu

3:1 et dixit Dominus ad me adhuc uade dilige mulierem 3:1 Then said the LORD unto me, Go yet, love a woman

beloved of her friend, yet an adulteress, according to the love of the LORD toward the children of Israel, who look

dilectam amico et adulteram sicut diligit Dominus filios Is- to other gods, and love flagons of wine. rahel et ipsi respectant ad deos alienos et diligunt uinacea uuarum 3:2 et fodi eam mihi quindecim argenteis et choro 3:2 So I bought her to me for fifteen pieces of silver, and

for an homer of barley, and an half homer of barley:

hordei et dimidio choro hordei 3:3 et dixi ad eam dies mul- 3:3 And I said unto her, Thou shalt abide for me many

days; thou shalt not play the harlot, and thou shalt not be for another man: so will I also be for thee.

tos expectabis me non fornicaberis et non eris uiro sed et ego expectabo te 3:4 quia dies multos sedebunt filii Israhel sine 3:4 For the children of Israel shall abide many days with-

out a king, and without a prince, and without a sacrifice, and without an image, and without an ephod, and without

rege et sine principe et sine sacrificio et sine altari et sine teraphim: ephod et sine therafin 3:5 et post haec reuertentur filii Isra- 3:5 Afterward shall the children of Israel return, and seek

the LORD their God, and David their king; and shall fear the LORD and his goodness in the latter days.

hel et quaerent Dominum Deum suum et Dauid regem suum et pauebunt ad Dominum et ad bonum eius in nouissimo dierum

4:1 audite uerbum Domini filii Israhel quia iudicium Do- 4:1 Hear the word of the LORD, ye children of Israel: for

the LORD hath a controversy with the inhabitants of the land, because there is no truth, nor mercy, nor knowledge

mino cum habitatoribus terrae non est enim ueritas et non est of God in the land. misericordia et non est scientia Dei in terra 4:2 maledictum 4:2 By swearing, and lying, and killing, and stealing, and

committing adultery, they break out, and blood toucheth blood.

et mendacium et homicidium et furtum et adulterium inundauerunt et sanguis sanguinem tetigit 4:3 propter hoc luge- 4:3 Therefore shall the land mourn, and every one that

dwelleth therein shall languish, with the beasts of the field, and with the fowls of heaven; yea, the fishes of the sea also

bit terra et infirmabitur omnis qui habitat in ea in bestia agri shall be taken away. et in uolucre caeli sed et pisces maris congregabuntur 4:4 4:4 Yet let no man strive, nor reprove another: for thy peo-

ple are as they that strive with the priest.

uerumtamen unusquisque non iudicet et non arguatur uir populus enim tuus sicut hii qui contradicunt sacerdoti 4:5 et 4:5 Therefore shalt thou fall in the day, and the prophet

also shall fall with thee in the night, and I will destroy thy mother.

corrues hodie et corruet etiam propheta tecum nocte tacere feci matrem tuam 4:6 conticuit populus meus eo quod non 4:6 My people are destroyed for lack of knowledge: be-

cause thou hast rejected knowledge, I will also reject thee, that thou shalt be no priest to me: seeing thou hast forgotten the law of thy God, I will also forget thy children.

Biblia Sacra Vulgata

habuerit scientiam quia tu scientiam reppulisti repellam te ne sacerdotio fungaris mihi et oblita es legis Dei tui obliuiscar

4:7 As they were increased, so they sinned against me: filiorum tuorum et ego 4:7 secundum multitudinem eorum

therefore will I change their glory into shame.

sic peccauerunt mihi gloriam eorum in ignominiam commu-

4:8 They eat up the sin of my people, and they set their tabo 4:8 peccata populi mei comedent et ad iniquitatem eo-

heart on their iniquity. 4:9 And there shall be, like people, like priest: and I will rum subleuabunt animas eorum 4:9 et erit sicut populus sic punish them for their ways, and reward them their doings.

sacerdos et uisitabo super eum uias eius et cogitationes eius

4:10 For they shall eat, and not have enough: they shall reddam ei 4:10 et comedent et non saturabuntur fornicati

commit whoredom, and shall not increase: because they have left off to take heed to the LORD.

sunt et non cessauerunt quoniam Dominum reliquerunt in

4:11 Whoredom and wine and new wine take away the non custodiendo 4:11 fornicatio et uinum et ebrietas aufert

heart.

4:12 My people ask counsel at their stocks, and their staff cor 4:12 populus meus in ligno suo interrogauit et baculus

declareth unto them: for the spirit of whoredoms hath caused them to err, and they have gone a whoring from

under their God. eius adnuntiauit ei spiritus enim fornicationum decepit eos et 4:13 They sacrifice upon the tops of the mountains, and fornicati sunt a Deo suo 4:13 super capita montium sacri-

burn incense upon the hills, under oaks and poplars and elms, because the shadow thereof is good: therefore your daughters shall commit whoredom, and your spouses shall commit adultery.

ficabant et super colles accendebant thymiama subtus quercum et populum et terebinthum quia bona erat umbra eius ideo fornicabuntur filiae uestrae et sponsae uestrae adulterae

4:14 I will not punish your daughters when they commit erunt 4:14 non uisitabo super filias uestras cum fuerint for-

whoredom, nor your spouses when they commit adultery: for themselves are separated with whores, and they sacrifice with harlots: therefore the people that doth not understand shall fall.

nicatae et super sponsas uestras cum adulterauerint quoniam ipsi cum meretricibus uersabantur et cum effeminatis sacri-

4:15 Though thou, Israel, play the harlot, yet let not Judah ficabant et populus non intellegens uapulabit 4:15 si forni-

offend; and come not ye unto Gilgal, neither go ye up to Bethaven, nor swear, The LORD liveth.

caris tu Israhel non delinquat saltim Iuda et nolite ingredi in Galgala et ne ascenderitis in Bethauen neque iuraueritis

4:16 For Israel slideth back as a backsliding heifer: now uiuit Dominus 4:16 quoniam sicut uacca lasciuiens decli-

the LORD will feed them as a lamb in a large place.

nauit Israhel nunc pascet eos Dominus quasi agnum in lati-

4:17 Ephraim is joined to idols: let him alone. tudine 4:17 particeps idolorum Ephraim dimitte eum 4:18

4:18 Their drink is sour: they have committed whoredom continually: her rulers with shame do love, Give ye.

separatum est conuiuium eorum fornicatione fornicati sunt

4:19 The wind hath bound her up in her wings, and they dilexerunt adferre ignominiam protectores eius 4:19 ligauit

shall be ashamed because of their sacrifices.

spiritus eam in alis suis et confundentur a sacrificiis suis

5:1 Hear ye this, O priests; and hearken, ye house of Is- 5:1 audite hoc sacerdotes et adtendite domus Israhel et do-

rael; and give ye ear, O house of the king; for judgment is toward you, because ye have been a snare on Mizpah, and

a net spread upon Tabor. mus regis auscultate quia uobis iudicium est quoniam la-

queus facti estis speculationi et rete expansum super Tha-

5:2 And the revolters are profound to make slaughter, bor 5:2 et uictimas declinastis in profundum et ego erudi-

though I have been a rebuker of them all.

5:3 I know Ephraim, and Israel is not hid from me: for tor omnium eorum 5:3 ego scio Ephraim et Israhel non est

now, O Ephraim, thou committest whoredom, and Israel is defiled.

Prophetia Osee

absconditus a me quia nunc fornicatus est Ephraim contaminatus est Israhel 5:4 non dabunt cogitationes suas ut re- 5:4 They will not frame their doings to turn unto their God:

for the spirit of whoredoms is in the midst of them, and they have not known the LORD.

uertantur ad Dominum suum quia spiritus fornicationis in

medio eorum et Dominum non cognouerunt 5:5 et respon- 5:5 And the pride of Israel doth testify to his face: therefore shall Israel and Ephraim fall in their iniquity: Judah also shall fall with them.

debit arrogantia Israhel in facie eius et Israhel et Ephraim ruent in iniquitate sua ruet etiam Iudas cum eis 5:6 in gre- 5:6 They shall go with their flocks and with their herds

to seek the LORD; but they shall not find him; he hath withdrawn himself from them.

gibus suis et in armentis suis uadent ad quaerendum Do-

minum et non inuenient ablatus est ab eis 5:7 in Domino 5:7 They have dealt treacherously against the LORD: for they have begotten strange children: now shall a month devour them with their portions.

praeuaricati sunt quia filios alienos genuerunt nunc deuorabit eos mensis cum partibus suis 5:8 clangite bucina in 5:8 Blow ye the cornet in Gibeah, and the trumpet in

Ramah: cry aloud at Bethaven, after thee, O Benjamin.

Gabaa tuba in Rama ululate in Bethauen post tergum tuum Beniamin 5:9 Ephraim in desolatione erit in die correptio- 5:9 Ephraim shall be desolate in the day of rebuke: among

the tribes of Israel have I made known that which shall surely be.

nis in tribubus Israhel ostendi fidem 5:10 facti sunt prin-

5:10 The princes of Judah were like them that remove the bound: therefore I will pour out my wrath upon them like water.

cipes Iuda quasi adsumentes terminum super eos effundam quasi aquam iram meam 5:11 calumniam patiens Ephraim 5:11 Ephraim is oppressed and broken in judgment, be-

cause he willingly walked after the commandment.

fractus iudicio quoniam coepit abire post sordem 5:12 et 5:12 Therefore will I be unto Ephraim as a moth, and to

the house of Judah as rottenness.

ego quasi tinea Ephraim et quasi putredo domui Iuda 5:13 5:13 When Ephraim saw his sickness, and Judah saw his

wound, then went Ephraim to the Assyrian, and sent to king Jareb: yet could he not heal you, nor cure you of

et uidit Ephraim languorem suum et Iudas uinculum suum your wound. et abiit Ephraim ad Assur et misit ad regem ultorem et ipse non poterit sanare uos nec soluere poterit a uobis uinculum 5:14 quoniam ego quasi leaena Ephraim et quasi catulus le- 5:14 For I will be unto Ephraim as a lion, and as a young

lion to the house of Judah: I, even I, will tear and go away; I will take away, and none shall rescue him.

onis domui Iuda ego ego capiam et uadam tollam et non est

qui eruat 5:15 uadens reuertar ad locum meum donec defi- 5:15 I will go and return to my place, till they acknowledge their offence, and seek my face: in their affliction they will seek me early.

ciatis et quaeratis faciem meam

6:1 in tribulatione sua mane consurgunt ad me uenite et 6:1 Come, and let us return unto the LORD: for he hath

torn, and he will heal us; he hath smitten, and he will bind us up.

reuertamur ad Dominum 6:2 quia ipse cepit et sanabit nos

6:2 After two days will he revive us: in the third day he will raise us up, and we shall live in his sight.

percutiet et curabit nos 6:3 uiuificabit nos post duos dies in _{6:3 Then shall we know, if we follow on to know the}

LORD: his going forth is prepared as the morning; and he shall come unto us as the rain, as the latter and former

die tertia suscitabit nos et uiuemus in conspectu eius scie- _{rain unto the earth.} mus sequemurque ut cognoscamus Dominum quasi diluculum praeparatus est egressus eius et ueniet quasi imber nobis temporaneus et serotinus terrae 6:4 quid faciam tibi Eph- 6:4 O Ephraim, what shall I do unto thee? O Judah, what

shall I do unto thee? for your goodness is as a morning cloud, and as the early dew it goeth away.

raim quid faciam tibi Iuda misericordia uestra quasi nubes matutina et quasi ros mane pertransiens 6:5 propter hoc 6:5 Therefore have I hewed them by the prophets; I have

slain them by the words of my mouth: and thy judgments are as the light that goeth forth.

Biblia Sacra Vulgata

dolaui in prophetis occidi eos in uerbis oris mei et iudi-

6:6 For I desired mercy, and not sacrifice; and the knowl- cia tua quasi lux egredientur 6:6 quia misericordiam uolui

edge of God more than burnt offerings.

et non sacrificium et scientiam Dei plus quam holocausta

6:7 But they like men have transgressed the covenant: 6:7 ipsi autem sicut Adam transgressi sunt pactum ibi prae-

there have they dealt treacherously against me. 6:8 Gilead is a city of them that work iniquity, and is pol- uaricati sunt in me 6:8 Galaad ciuitas operantium idolum luted with blood.

6:9 And as troops of robbers wait for a man, so the com- subplantata sanguine 6:9 et quasi fauces uirorum latronum

pany of priests murder in the way by consent: for they commit lewdness.

particeps sacerdotum in uia interficientium pergentes de Sy-

6:10 I have seen an horrible thing in the house of Israel: chem quia scelus operati sunt 6:10 in domo Israhel uidi hor-

there is the whoredom of Ephraim, Israel is defiled.

rendum ibi fornicationes Ephraim contaminatus est Israhel

6:11 Also, O Judah, he hath set an harvest for thee, when 6:11 sed et Iuda pone messem tibi cum conuertero captiuita-

I returned the captivity of my people.

tem populi mei

7:1 When I would have healed Israel, then the iniquity of 7:1 cum sanare uellem Israhel reuelata est iniquitas Eph-

Ephraim was discovered, and the wickedness of Samaria: for they commit falsehood; and the thief cometh in, and

the troop of robbers spoileth without. raim et malitia Samariae quia operati sunt mendacium et 7:2 And they consider not in their hearts that I remem- fur ingressus est spolians latrunculus foris 7:2 et ne forte

ber all their wickedness: now their own doings have beset them about; they are before my face.

dicant in cordibus suis omnem malitiam eorum me recordatum nunc circumdederunt eos adinuentiones suae coram

7:3 They make the king glad with their wickedness, and facie mea factae sunt 7:3 in malitia sua laetificauerunt re-

the princes with their lies. 7:4 They are all adulterers, as an oven heated by the baker, gem et in mendaciis suis principes 7:4 omnes adulterantes who ceaseth from raising after he hath kneaded the dough, until it be leavened.

quasi clibanus succensus a coquente quieuit paululum ciui-

7:5 In the day of our king the princes have made him tas a commixtione fermenti donec fermentaretur totum 7:5

sick with bottles of wine; he stretched out his hand with scorners.

dies regis nostri coeperunt principes furere a uino extendit

7:6 For they have made ready their heart like an oven, manum suam cum inlusoribus 7:6 quia adplicuerunt quasi

whiles they lie in wait: their baker sleepeth all the night; in the morning it burneth as a flaming fire.

clibanum cor suum cum insidiaretur eis tota nocte dormiuit

7:7 They are all hot as an oven, and have devoured their coquens eos mane ipse succensus quasi ignis flammae 7:7

judges; all their kings are fallen: there is none among them that calleth unto me.

omnes calefacti sunt quasi clibanus et deuorauerunt iudices suos omnes reges eorum ceciderunt non est qui clamet in eis

7:8 Ephraim, he hath mixed himself among the people; ad me 7:8 Ephraim in populis ipse commiscebatur Ephraim

Ephraim is a cake not turned.

7:9 Strangers have devoured his strength, and he knoweth factus est subcinericius qui non reuersatur 7:9 comederunt

it not: yea, gray hairs are here and there upon him, yet he knoweth not.

alieni robur eius et ipse nesciuit sed et cani effusi sunt in

7:10 And the pride of Israel testifieth to his face: and they eo et ipse ignorauit 7:10 et humiliabitur superbia Israhel in

do not return to the LORD their God, nor seek him for all this.

facie eius nec reuersi sunt ad Dominum Deum suum et non

7:11 Ephraim also is like a silly dove without heart: they quaesierunt eum in omnibus his 7:11 et factus est Ephraim call to Egypt, they go to Assyria.

quasi columba seducta non habens cor Aegyptum inuoca-

Prophetia Osee

bant ad Assyrios abierunt 7:12 et cum profecti fuerint ex- 7:12 When they shall go, I will spread my net upon them;

I will bring them down as the fowls of the heaven; I will chastise them, as their congregation hath heard.

pandam super eos rete meum quasi uolucrem caeli detraham eos caedam eos secundum auditionem coetus eorum 7:13 7:13 Woe unto them! for they have fled from me: destruc-

tion unto them! because they have transgressed against me: though I have redeemed them, yet they have spoken

uae eis quoniam recesserunt a me uastabuntur quia praeuari- lies against me. cati sunt in me et ego redeei eos et ipsi locuti sunt contra me mendacia 7:14 et non clamauerunt ad me in corde suo sed 7:14 And they have not cried unto me with their heart,

when they howled upon their beds: they assemble themselves for corn and wine, and they rebel against me.

ululabant in cubilibus suis super triticum et uinum ruminabant recesserunt a me 7:15 et ego erudiui et confortaui bra- 7:15 Though I have bound and strengthened their arms,

yet do they imagine mischief against me.

chia eorum et in me cogitauerunt malitiam 7:16 reuersi sunt 7:16 They return, but not to the most High: they are like a

deceitful bow: their princes shall fall by the sword for the rage of their tongue: this shall be their derision in the land

ut essent absque iugo facti sunt quasi arcus dolosus cadent of Egypt. in gladio principes eorum a furore linguae suae ista subsannatio eorum in terra Aegypti

8:1 in gutture tuo sit tuba quasi aquila super domum Do- 8:1 Set the trumpet to thy mouth. He shall come as an

eagle against the house of the LORD, because they have transgressed my covenant, and trespassed against my law.

mini pro eo quod transgressi sunt foedus meum et legem meam praeuaricati sunt 8:2 me inuocabunt Deus meus co- 8:2 Israel shall cry unto me, My God, we know thee. gnouimus te Israhel 8:3 proiecit Israhel bonum inimicus 8:3 Israel hath cast off the thing that is good: the enemy

shall pursue him.

persequetur eum 8:4 ipsi regnauerunt et non ex me prin- 8:4 They have set up kings, but not by me: they have made

princes, and I knew it not: of their silver and their gold have they made them idols, that they may be cut off.

cipes extiterunt et non cognoui argentum suum et aurum suum fecerunt sibi idola ut interirent 8:5 proiectus est ui- 8:5 Thy calf, O Samaria, hath cast thee off; mine anger is

kindled against them: how long will it be ere they attain to innocency?

tulus tuus Samaria iratus est furor meus in eis usquequo non poterunt emundari 8:6 quia ex Israhel et ipse est artifex fe- 8:6 For from Israel was it also: the workman made it;

therefore it is not God: but the calf of Samaria shall be broken in pieces.

cit illum et non est Deus quoniam in aranearum telas erit uitulus Samariae 8:7 quia uentum seminabunt et turbinem 8:7 For they have sown the wind, and they shall reap the

whirlwind: it hath no stalk; the bud shall yield no meal: if so be it yield, the strangers shall swallow it up.

metent culmus stans non est in eis germen non faciet farinam quod si et fecerit alieni comedent eam 8:8 deuoratus 8:8 Israel is swallowed up: now shall they be among the

Gentiles as a vessel wherein is no pleasure.

est Israhel nunc factus est in nationibus quasi uas inmundum 8:9 quia ipsi ascenderunt ad Assur onager solitarius 8:9 For they are gone up to Assyria, a wild ass alone by

himself: Ephraim hath hired lovers.

sibi Ephraim munera dederunt amatoribus 8:10 sed et cum 8:10 Yea, though they have hired among the nations, now

will I gather them, and they shall sorrow a little for the burden of the king of princes.

mercede conduxerint nationes nunc congregabo eos et quiescent paulisper ab onere regis et principum 8:11 quia multi- 8:11 Because Ephraim hath made many altars to sin, altars

shall be unto him to sin.

plicauit Ephraim altaria ad peccandum factae sunt ei arae in delictum 8:12 scribam ei multiplices leges meas quae ue- 8:12 I have written to him the great things of my law, but

they were counted as a strange thing.

lut alienae conputatae sunt 8:13 hostias adfer adfer immo- 8:13 They sacrifice flesh for the sacrifices of mine offer-

ings, and eat it; but the LORD accepteth them not; now will he remember their iniquity, and visit their sins: they shall return to Egypt.

Biblia Sacra Vulgata

labunt carnes et comedent Dominus non suscipiet eas nunc recordabitur iniquitatis eorum et uisitabit peccata eorum ipsi

8:14 For Israel hath forgotten his Maker, and buildeth tem- in Aegyptum conuertentur 8:14 et oblitus est Israhel facples; and Judah hath multiplied fenced cities: but I will send a fire upon his cities, and it shall devour the palaces

thereof. toris sui et aedificauit delubra et Iudas multiplicauit urbes munitas et mittam ignem in ciuitates eius et deuorabit aedes illius

9:1 Rejoice not, O Israel, for joy, as other people: for thou 9:1 noli laetari Israhel noli exultare sicut populi quia for-

hast gone a whoring from thy God, thou hast loved a reward upon every cornfloor.

nicatus es a Deo tuo dilexisti mercedem super omnes areas

9:2 The floor and the winepress shall not feed them, and tritici 9:2 area et torcular non pascet eos et uinum mentie-

the new wine shall fail in her.

9:3 They shall not dwell in the LORD’s land; but Ephraim tur eis 9:3 non habitabunt in terra Domini reuersus est Eph-

shall return to Egypt, and they shall eat unclean things in Assyria.

raim Aegyptum et in Assyriis pollutum comedit 9:4 non li-

9:4 They shall not offer wine offerings to the LORD, neither shall they be pleasing unto him: their sacrifices shall be unto them as the bread of mourners; all that eat thereof shall be polluted: for their bread for their soul shall not come into the house of the LORD.

babunt Domino uinum et non placebunt ei sacrificia eorum quasi panis lugentium omnes qui comedunt eum contaminabuntur quia panis eorum animae ipsorum non intrabit in

9:5 What will ye do in the solemn day, and in the day of domum Domini 9:5 quid facietis in die sollemni in die fes-

the feast of the LORD?

9:6 For, lo, they are gone because of destruction: Egypt tiuitatis Domini 9:6 ecce enim profecti sunt a uastitate Ae-

shall gather them up, Memphis shall bury them: the pleasant places for their silver, nettles shall possess them:

thorns shall be in their tabernacles. gyptus congregauit eos Memphis sepeliet eos desiderabile

argenti eorum urtica hereditabit lappa in tabernaculis eorum

9:7 The days of visitation are come, the days of recom- 9:7 uenerunt dies uisitationis uenerunt dies retributionis sci-

pence are come; Israel shall know it: the prophet is a fool, the spiritual man is mad, for the multitude of thine iniq-

uity, and the great hatred. tote Israhel stultum prophetam insanum uirum spiritalem

propter multitudinem iniquitatis tuae et multitudo amentiae

9:8 The watchman of Ephraim was with my God: but the 9:8 speculator Ephraim cum Deo meo propheta laqueus rui-

prophet is a snare of a fowler in all his ways, and hatred in the house of his God.

nae super omnes uias eius insania in domo Dei eius 9:9

9:9 They have deeply corrupted themselves, as in the days of Gibeah: therefore he will remember their iniquity, he will visit their sins.

profunde peccauerunt sicut in diebus Gabaa recordabitur in-

9:10 I found Israel like grapes in the wilderness; I saw iquitatis eorum et uisitabit peccata eorum 9:10 quasi uuas

your fathers as the firstripe in the fig tree at her first time: but they went to Baalpeor, and separated themselves unto that shame; and their abominations were according as they loved.

in deserto inueni Israhel quasi prima poma ficulneae in cacumine eius uidi patres eorum ipsi autem intrauerunt ad Beelphegor et abalienati sunt in confusione et facti sunt abo-

9:11 As for Ephraim, their glory shall fly away like a bird, minabiles sicut ea quae dilexerunt 9:11 Ephraim quasi auis

from the birth, and from the womb, and from the conception.

auolauit gloria eorum a partu et ab utero et a conceptu 9:12

9:12 Though they bring up their children, yet will I bereave them, that there shall not be a man left: yea, woe also to them when I depart from them!

quod si et enutrierint filios suos absque liberis eos faciam in

9:13 Ephraim, as I saw Tyrus, is planted in a pleasant hominibus sed et uae eis cum recessero ab eis 9:13 Ephraim

place: but Ephraim shall bring forth his children to the murderer.

ut uidi Tyrus erat fundata in pulchritudine et Ephraim educit

Prophetia Osee

ad interfectorem filios suos 9:14 da eis Domine quid dabis 9:14 Give them, O LORD: what wilt thou give? give them

a miscarrying womb and dry breasts.

eis da eis uuluam sine liberis et ubera arentia 9:15 omnes 9:15 All their wickedness is in Gilgal: for there I hated

them: for the wickedness of their doings I will drive them out of mine house, I will love them no more: all their

nequitiae eorum in Galgal quia ibi exosos habui eos prop- princes are revolters. ter malitiam adinuentionum eorum de domo mea eiciam eos non addam ut diligam eos omnes principes eorum receden-

tes 9:16 percussus est Ephraim radix eorum exsiccata est 9:16 Ephraim is smitten, their root is dried up, they shall bear no fruit: yea, though they bring forth, yet will I slay even the beloved fruit of their womb.

fructum nequaquam facient quod si et genuerint interficiam amantissima uteri eorum 9:17 abiciet eos Deus meus quia 9:17 My God will cast them away, because they did not

hearken unto him: and they shall be wanderers among the nations.

non audierunt eum et erunt uagi in nationibus

10:1 uitis frondosa Israhel fructus adaequatus est ei se- 10:1 Israel is an empty vine, he bringeth forth fruit unto

himself: according to the multitude of his fruit he hath increased the altars; according to the goodness of his land

cundum multitudinem fructus sui multiplicauit altaria iuxta they have made goodly images. ubertatem terrae suae exuberauit simulacris 10:2 diuisum 10:2 Their heart is divided; now shall they be found faulty:

he shall break down their altars, he shall spoil their images.

est cor eorum nunc interibunt ipse confringet simulacra eorum depopulabitur aras eorum 10:3 quia nunc dicent non 10:3 For now they shall say, We have no king, because we

feared not the LORD; what then should a king do to us?

est rex nobis non enim timemus Dominum et rex quid faciet nobis 10:4 loquimini uerba uisionis inutilis et ferietis 10:4 They have spoken words, swearing falsely in making

a covenant: thus judgment springeth up as hemlock in the furrows of the field.

foedus et germinabit quasi amaritudo iudicium super sulcos agri 10:5 uaccas Bethauen coluerunt habitatores Samariae 10:5 The inhabitants of Samaria shall fear because of the

calves of Bethaven: for the people thereof shall mourn over it, and the priests thereof that rejoiced on it, for the

quia luxit super eum populus eius et aeditui eius super eum glory thereof, because it is departed from it. exultauerunt in gloria eius quia migrauit ab eo 10:6 siqui- 10:6 It shall be also carried unto Assyria for a present to

king Jareb: Ephraim shall receive shame, and Israel shall be ashamed of his own counsel.

dem et ipse in Assur delatus est munus regi ultori confusio Ephraim capiet et confundetur Israhel in uoluntate sua 10:7 10:7 As for Samaria, her king is cut off as the foam upon

the water.

transire fecit Samaria regem suum quasi spumam super faciem aquae 10:8 et disperdentur excelsa idoli peccatum Is- 10:8 The high places also of Aven, the sin of Israel, shall

be destroyed: the thorn and the thistle shall come up on their altars; and they shall say to the mountains, Cover us;

rahel lappa et tribulus ascendet super aras eorum et dicent and to the hills, Fall on us. montibus operite nos et collibus cadite super nos 10:9 ex 10:9 O Israel, thou hast sinned from the days of Gibeah:

there they stood: the battle in Gibeah against the children of iniquity did not overtake them.

diebus Gabaa peccauit Israhel ibi steterunt non conprehendet eos in Gabaa proelium super filios iniquitatis 10:10 iuxta 10:10 It is in my desire that I should chastise them; and

the people shall be gathered against them, when they shall bind themselves in their two furrows.

desiderium meum corripiam eos congregabuntur super eos populi cum corripientur propter duas iniquitates suas 10:11 10:11 And Ephraim is as an heifer that is taught, and

loveth to tread out the corn; but I passed over upon her fair neck: I will make Ephraim to ride; Judah shall plow,

Ephraim uitula docta diligere trituram et ego transiui su- and Jacob shall break his clods. per pulchritudinem colli eius ascendam super Ephraim arabit Iudas confringet sibi sulcos Iacob 10:12 seminate uobis 10:12 Sow to yourselves in righteousness, reap in mercy;

break up your fallow ground: for it is time to seek the LORD, till he come and rain righteousness upon you.

Biblia Sacra Vulgata

in iustitia metite in ore misericordiae innouate uobis nouale tempus autem requirendi Dominum cum uenerit qui docebit

10:13 Ye have plowed wickedness, ye have reaped iniq- uos iustitiam 10:13 arastis impietatem iniquitatem messui-

uity; ye have eaten the fruit of lies: because thou didst trust in thy way, in the multitude of thy mighty men.

stis comedistis frugem mendacii quia confisus es in uiis tuis

10:14 Therefore shall a tumult arise among thy people, in multitudine fortium tuorum 10:14 consurget tumultus in

and all thy fortresses shall be spoiled, as Shalman spoiled Betharbel in the day of battle: the mother was dashed in

pieces upon her children. populo tuo et omnes munitiones tuae uastabuntur sicut uas-

tatus est Salman a domo eius qui iudicauit Baal in die proelii

10:15 So shall Bethel do unto you because of your great matre super filios adlisa 10:15 sic fecit uobis Bethel a facie

wickedness: in a morning shall the king of Israel utterly be cut off.

malitiae nequitiarum uestrarum

11:1 When Israel was a child, then I loved him, and called 11:1 sicuti mane transit pertransiit rex Israhel quia puer Is-

my son out of Egypt.

11:2 As they called them, so they went from them: they rahel et dilexi eum et ex Aegypto uocaui filium meum 11:2

sacrificed unto Baalim, and burned incense to graven images.

uocauerunt eos sic abierunt a facie eorum Baalim immola-

11:3 I taught Ephraim also to go, taking them by their bant et simulacris sacrificabant 11:3 et ego quasi nutricius

arms; but they knew not that I healed them.

Ephraim portabam eos in brachiis meis et nescierunt quod

11:4 I drew them with cords of a man, with bands of love: curarem eos 11:4 in funiculis Adam traham eos in uinculis

and I was to them as they that take off the yoke on their jaws, and I laid meat unto them.

caritatis et ero eis quasi exaltans iugum super maxillas eo-

11:5 He shall not return into the land of Egypt, and the rum et declinaui ad eum ut uesceretur 11:5 non reuertetur

Assyrian shall be his king, because they refused to return.

in terram Aegypti et Assur ipse rex eius quoniam noluerunt

11:6 And the sword shall abide on his cities, and shall con- conuerti 11:6 coepit gladius in ciuitatibus eius et consumet

sume his branches, and devour them, because of their own counsels.

electos eius et comedet capita eorum 11:7 et populus meus

11:7 And my people are bent to backsliding from me: though they called them to the most High, none at all would exalt him.

pendebit ad rediium meum iugum autem inponetur ei simul

11:8 How shall I give thee up, Ephraim? how shall I de- quod non auferetur 11:8 quomodo dabo te Ephraim prote-

liver thee, Israel? how shall I make thee as Admah? how shall I set thee as Zeboim? mine heart is turned within me,

my repentings are kindled together. gam te Israhel quomodo dabo te sicut Adama ponam te ut

Seboim conuersum est in me cor meum pariter conturbata

11:9 I will not execute the fierceness of mine anger, I will est paenitudo mea 11:9 non faciam furorem irae meae non not return to destroy Ephraim: for I am God, and not man; the Holy One in the midst of thee: and I will not enter into

the city. conuertar ut disperdam Ephraim quoniam Deus ego et non 11:10 They shall walk after the LORD: he shall roar like homo in medio tui Sanctus et non ingrediar ciuitatem 11:10

a lion: when he shall roar, then the children shall tremble from the west.

post Dominum ambulabunt quasi leo rugiet quia ipse rugiet

11:11 They shall tremble as a bird out of Egypt, and as a et formidabunt filii maris 11:11 et auolabunt quasi auis ex

dove out of the land of Assyria: and I will place them in their houses, saith the LORD.

Aegypto et quasi columba de terra Assyriorum et conlocabo

11:12 Ephraim compasseth me about with lies, and the eos in domibus suis dicit Dominus 11:12 circumdedit me

house of Israel with deceit: but Judah yet ruleth with God, and is faithful with the saints.

in negatione Ephraim et in dolo domus Israhel Iudas autem testis descendit cum Deo et cum sanctis fidelis

Prophetia Osee

12:1 Ephraim pascit uentum et sequitur aestum tota die 12:1 Ephraim feedeth on wind, and followeth after the east wind: he daily increaseth lies and desolation; and they do make a covenant with the Assyrians, and oil is carried into

mendacium et uastitatem multiplicat et foedus cum Assyriis Egypt. iniit et oleum in Aegyptum ferebat 12:2 iudicium ergo Do- 12:2 The LORD hath also a controversy with Judah, and

will punish Jacob according to his ways; according to his doings will he recompense him.

mini cum Iuda et uisitatio super Iacob iuxta uias eius et iuxta adinuentiones eius reddet ei 12:3 in utero subplantauit frat- 12:3 He took his brother by the heel in the womb, and by

his strength he had power with God:

rem suum et in fortitudine sua directus est cum angelo 12:4 12:4 Yea, he had power over the angel, and prevailed: he

wept, and made supplication unto him: he found him in Bethel, and there he spake with us;

et inualuit ad angelum et confortatus est fleuit et rogauit eum in Bethel inuenit eum et ibi locutus est nobiscum 12:5 et 12:5 Even the LORD God of hosts; the LORD is his

memorial.

Dominus Deus exercituum Dominus memoriale eius 12:6 12:6 Therefore turn thou to thy God: keep mercy and judg-

ment and wait on thy God continually.

et tu ad Deum tuum conuerteris misericordiam et iudicium custodi et spera in Deo tuo semper 12:7 Chanaan in manu 12:7 He is a merchant, the balances of deceit are in his

hand: he loveth to oppress.

eius statera dolosa calumniam dilexit 12:8 et dixit Ephraim 12:8 And Ephraim said, Yet I am become rich, I have

found me out substance: in all my labours they shall find none iniquity in me that were sin.

uerumtamen diues effectus sum inueni idolum mihi omnes labores mei non inuenient mihi iniquitatem quam peccaui 12:9 et ego Dominus Deus tuus ex terra Aegypti adhuc sedere 12:9 And I that am the LORD thy God from the land of

Egypt will yet make thee to dwell in tabernacles, as in the days of the solemn feast.

te faciam in tabernaculis sicut in diebus festiuitatis 12:10 et

12:10 I have also spoken by the prophets, and I have multiplied visions, and used similitudes, by the ministry of the prophets.

locutus sum super prophetas et ego uisionem multiplicaui

et in manu prophetarum adsimilatus sum 12:11 si Galaad 12:11 Is there iniquity in Gilead? surely they are vanity: they sacrifice bullocks in Gilgal; yea, their altars are as heaps in the furrows of the fields.

idolum tamen frustra erant in Galgal bubus immolantes nam et altaria eorum quasi acerui super sulcos agri 12:12 fugit 12:12 And Jacob fled into the country of Syria, and Israel

served for a wife, and for a wife he kept sheep.

Iacob in regionem Syriae et seruiuit Israhel in uxore et in uxore seruauit 12:13 in propheta autem eduxit Dominus Is- 12:13 And by a prophet the LORD brought Israel out of

Egypt, and by a prophet was he preserved.

rahel de Aegypto et in propheta seruatus est 12:14 ad ira- 12:14 Ephraim provoked him to anger most bitterly: therefore shall he leave his blood upon him, and his reproach shall his LORD return unto him.

cundiam me prouocauit Ephraim in amaritudinibus suis et sanguis eius super eum ueniet et obprobrium eius restituet ei Dominus suus

13:1 loquente Ephraim horror inuasit Israhel et deliquit in 13:1 When Ephraim spake trembling, he exalted himself

in Israel; but when he offended in Baal, he died.

Baal et mortuus est 13:2 et nunc addiderunt ad peccandum 13:2 And now they sin more and more, and have made

them molten images of their silver, and idols according to their own understanding, all of it the work of the craftsmen: they say of them, Let the men that sacrifice kiss the calves.

feceruntque sibi conflatile de argento suo quasi similitudinem idolorum factura artificum totum est his ipsi dicunt immolate homines uitulos adorantes 13:3 idcirco erunt quasi 13:3 Therefore they shall be as the morning cloud and as

the early dew that passeth away, as the chaff that is driven with the whirlwind out of the floor, and as the smoke out

nubes matutina et sicut ros matutinus praeteriens sicut puluis of the chimney. turbine raptus ex area et sicut fumus de fumario 13:4 ego 13:4 Yet I am the LORD thy God from the land of Egypt,

and thou shalt know no god but me: for there is no saviour beside me.

Biblia Sacra Vulgata

autem Dominus Deus tuus ex terra Aegypti et Deum absque

13:5 I did know thee in the wilderness, in the land of great me nescies et saluator non est praeter me 13:5 ego cognoui drought.

13:6 According to their pasture, so were they filled; they te in deserto in terra solitudinis 13:6 iuxta pascua sua et ad-

were filled, and their heart was exalted; therefore have they forgotten me.

impleti sunt et saturati eleuauerunt cor suum et obliti sunt

13:7 Therefore I will be unto them as a lion: as a leopard mei 13:7 et ero eis quasi leaena sicut pardus in uia Assy-

by the way will I observe them:

13:8 I will meet them as a bear that is bereaved of her riorum 13:8 occurram eis quasi ursa raptis catulis et dis-

whelps, and will rend the caul of their heart, and there will I devour them like a lion: the wild beast shall tear them.

rumpam interiora iecoris eorum et consumam eos ibi quasi

13:9 O Israel, thou hast destroyed thyself; but in me is leo bestia agri scindet eos 13:9 perditio tua Israhel tantum-

thine help. 13:10 I will be thy king: where is any other that may save modo in me auxilium tuum 13:10 ubi est rex tuus maxime thee in all thy cities? and thy judges of whom thou saidst, Give me a king and princes?

nunc saluet te in omnibus urbibus tuis et iudices tui de quibus

13:11 I gave thee a king in mine anger, and took him away dixisti da mihi regem et principes 13:11 dabo tibi regem in

in my wrath.

13:12 The iniquity of Ephraim is bound up; his sin is hid. furore meo et auferam in indignatione mea 13:12 conligata 13:13 The sorrows of a travailing woman shall come upon est iniquitas Ephraim absconditum peccatum eius 13:13 do-

him: he is an unwise son; for he should not stay long in the place of the breaking forth of children.

lores parturientis uenient ei ipse filius non sapiens nunc enim

13:14 I will ransom them from the power of the grave; non stabit in contritione filiorum 13:14 de manu mortis li-

I will redeem them from death: O death, I will be thy plagues; O grave, I will be thy destruction: repentance

shall be hid from mine eyes. berabo eos de morte rediiam eos ero mors tua o mors ero

morsus tuus inferne consolatio abscondita est ab oculis meis

13:15 Though he be fruitful among his brethren, an east 13:15 quia ipse inter fratres diuidet adducet urentem uentum

wind shall come, the wind of the LORD shall come up from the wilderness, and his spring shall become dry, and his fountain shall be dried up: he shall spoil the treasure of all pleasant vessels.

Dominus de deserto ascendentem et siccabit uenas eius et desolabit fontem eius et ipse diripiet thesaurum omnis uasis desiderabilis

14:1 Samaria shall become desolate; for she hath rebelled 14:1 pereat Samaria quoniam ad amaritudinem concitauit against her God: they shall fall by the sword: their infants shall be dashed in pieces, and their women with child shall

be ripped up. Dominum suum in gladio pereat paruuli eorum elidantur et 14:2 O israel, return unto the LORD thy God; for thou hast fetae eius discindantur 14:2 conuertere Israhel ad Domi-

fallen by thine iniquity.

14:3 Take with you words, and turn to the LORD: say unto num Deum tuum quoniam corruisti in iniquitate tua 14:3

him, Take away all iniquity, and receive us graciously: so will we render the calves of our lips.

tollite uobiscum uerba et conuertimini ad Dominum dicite ei omnem aufer iniquitatem et accipe bonum et reddemus uitu-

14:4 Asshur shall not save us; we will not ride upon los labiorum nostrorum 14:4 Assur non saluabit nos super

horses: neither will we say any more to the work of our hands, Ye are our gods: for in thee the fatherless findeth

mercy. equum non ascendemus nec dicemus ultra dii nostri opera

manuum nostrarum quia eius qui in te est misereberis pu-

14:5 I will heal their backsliding, I will love them freely: pilli 14:5 sanabo contritiones eorum diligam eos spontanee

for mine anger is turned away from him.

14:6 I will be as the dew unto Israel: he shall grow as the quia auersus est furor meus ab eo 14:6 ero quasi ros Israhel

lily, and cast forth his roots as Lebanon.

14:7 His branches shall spread, and his beauty shall be as germinabit quasi lilium et erumpet radix eius ut Libani 14:7

the olive tree, and his smell as Lebanon.

Prophetia Osee

ibunt rami eius et erit quasi oliua gloria eius et odor eius ut Libani 14:8 conuertentur sedentes in umbra eius uiuent tri- 14:8 They that dwell under his shadow shall return; they

shall revive as the corn, and grow as the vine: the scent thereof shall be as the wine of Lebanon.

tico et germinabunt quasi uinea memoriale eius sicut uinum Libani 14:9 Ephraim quid mihi ultra idola ego exaudiam et 14:9 Ephraim shall say, What have I to do any more with

idols? I have heard him, and observed him: I am like a green fir tree. From me is thy fruit found.

dirigam eum ego ut abietem uirentem ex me fructus tuus inuentus est 14:10 quis sapiens et intelleget ista intellegens et 14:10 Who is wise, and he shall understand these things?

prudent, and he shall know them? for the ways of the LORD are right, and the just shall walk in them: but the

sciet haec quia rectae uiae Domini et iusti ambulabunt in eis transgressors shall fall therein. praeuaricatores uero corruent in eis

Biblia Sacra Vulgata

MCCLXXIV

Prophetia Ioel

1:1 uerbum Domini quod factum est ad Iohel filium Fa- 1:1 The word of the LORD that came to Joel the son of

Pethuel.

tuhel 1:2 audite hoc senes et auribus percipite omnes ha- 1:2 Hear this, ye old men, and give ear, all ye inhabitants

of the land. Hath this been in your days, or even in the days of your fathers?

bitatores terrae si factum est istud in diebus uestris aut in diebus patrum uestrorum 1:3 super hoc filiis uestris narrate 1:3 Tell ye your children of it, and let your children tell

their children, and their children another generation.

et filii uestri filiis suis et filii eorum generationi alterae 1:4 1:4 That which the palmerworm hath left hath the locust

eaten; and that which the locust hath left hath the cankerworm eaten; and that which the cankerworm hath left hath

residuum erucae comedit lucusta et residuum lucustae co- the caterpiller eaten. medit bruchus et residuum bruchi comedit rubigo 1:5 ex- 1:5 Awake, ye drunkards, and weep; and howl, all ye

drinkers of wine, because of the new wine; for it is cut off from your mouth.

pergescimini ebrii et flete et ululate omnes qui bibitis uinum in dulcedine quoniam periit ab ore uestro 1:6 gens enim 1:6 For a nation is come up upon my land, strong, and

without number, whose teeth are the teeth of a lion, and he hath the cheek teeth of a great lion.

ascendit super terram meam fortis et innumerabilis dentes eius ut dentes leonis et molares eius ut catuli leonis 1:7 1:7 He hath laid my vine waste, and barked my fig tree:

he hath made it clean bare, and cast it away; the branches thereof are made white.

posuit uineam meam in desertum et ficum meam decorticauit nudans spoliauit eam et proiecit albi facti sunt rami eius 1:8 plange quasi uirgo accincta sacco super uirum pubertatis 1:8 Lament like a virgin girded with sackcloth for the hus-

band of her youth.

suae 1:9 periit sacrificium et libatio de domo Domini luxe- 1:9 The meat offering and the drink offering is cut off from

the house of the LORD; the priests, the LORD’s ministers, mourn.

runt sacerdotes ministri Domini 1:10 depopulata est regio

1:10 The field is wasted, the land mourneth; for the corn is wasted: the new wine is dried up, the oil languisheth.

luxit humus quoniam deuastatum est triticum confusum est uinum elanguit oleum 1:11 confusi sunt agricolae ululaue- 1:11 Be ye ashamed, O ye husbandmen; howl, O ye vine-

dressers, for the wheat and for the barley; because the harvest of the field is perished.

runt uinitores super frumento et hordeo quia periit messis agri 1:12 uinea confusa est et ficus elanguit malogranatum 1:12 The vine is dried up, and the fig tree languisheth; the

pomegranate tree, the palm tree also, and the apple tree, even all the trees of the field, are withered: because joy is

et palma et malum et omnia ligna agri aruerunt quia con- withered away from the sons of men. fusum est gaudium a filiis hominum 1:13 accingite uos et 1:13 Gird yourselves, and lament, ye priests: howl, ye

ministers of the altar: come, lie all night in sackcloth, ye ministers of my God: for the meat offering and the drink

plangite sacerdotes ululate ministri altaris ingredimini cu- offering is withholden from the house of your God. bate in sacco ministri Dei mei quoniam interiit de domo Dei uestri sacrificium et libatio 1:14 sanctificate ieiunium uo- 1:14 Sanctify ye a fast, call a solemn assembly, gather the

elders and all the inhabitants of the land into the house of the LORD your God, and cry unto the LORD,

cate coetum congregate senes omnes habitatores terrae in

domum Dei uestri et clamate ad Dominum 1:15 a a a diei 1:15 Alas for the day! for the day of the LORD is at hand,

and as a destruction from the Almighty shall it come.

quia prope est dies Domini et quasi uastitas a potente ueniet 1:16 numquid non coram oculis uestris alimenta perierunt de 1:16 Is not the meat cut off before our eyes, yea, joy and

gladness from the house of our God?

domo Dei nostri laetitia et exultatio 1:17 conputruerunt iu- 1:17 The seed is rotten under their clods, the garners are

laid desolate, the barns are broken down; for the corn is withered.

menta in stercore suo demolita sunt horrea dissipatae sunt apothecae quoniam confusum est triticum 1:18 quid inge- 1:18 How do the beasts groan! the herds of cattle are per-

plexed, because they have no pasture; yea, the flocks of sheep are made desolate.

muit animal mugierunt greges armenti quia non est pascua

Biblia Sacra Vulgata

1:19 O LORD, to thee will I cry: for the fire hath devoured eis sed et greges pecorum disperierunt 1:19 ad te Domine

the pastures of the wilderness, and the flame hath burned all the trees of the field.

clamabo quia ignis comedit speciosa deserti et flamma suc-

1:20 The beasts of the field cry also unto thee: for the cendit omnia ligna regionis 1:20 sed et bestiae agri quasi

rivers of waters are dried up, and the fire hath devoured the pastures of the wilderness.

area sitiens imbrem suspexerunt ad te quoniam exsiccati sunt fontes aquarum et ignis deuorauit speciosa deserti

2:1 Blow ye the trumpet in Zion, and sound an alarm in my 2:1 canite tuba in Sion ululate in monte sancto meo con-

holy mountain: let all the inhabitants of the land tremble: for the day of the LORD cometh, for it is nigh at hand;

turbentur omnes habitatores terrae quia uenit dies Domini

2:2 A day of darkness and of gloominess, a day of clouds quia prope est 2:2 dies tenebrarum et caliginis dies nubis et

and of thick darkness, as the morning spread upon the mountains: a great people and a strong; there hath not been ever the like, neither shall be any more after it, even to the years of many generations.

turbinis quasi mane expansum super montes populus multus et fortis similis ei non fuit a principio et post eum non erit

2:3 A fire devoureth before them; and behind them a flame usque in annos generationis et generationis 2:3 ante faciem

burneth: the land is as the garden of Eden before them, and behind them a desolate wilderness; yea, and nothing shall

escape them. eius ignis uorans et post eum exurens flamma quasi hortus

uoluptatis terra coram eo et post eum solitudo deserti neque

2:4 The appearance of them is as the appearance of horses; est qui effugiat eum 2:4 quasi aspectus equorum aspectus and as horsemen, so shall they run.

2:5 Like the noise of chariots on the tops of mountains eorum et quasi equites sic current 2:5 sicut sonitus quadri-

shall they leap, like the noise of a flame of fire that devoureth the stubble, as a strong people set in battle array.

garum super capita montium exilient sicut sonitus flammae ignis deuorantis stipulam uelut populus fortis praeparatus ad

2:6 Before their face the people shall be much pained: all proelium 2:6 a facie eius cruciabuntur populi omnes uultus

faces shall gather blackness.

2:7 They shall run like mighty men; they shall climb the rediientur in ollam 2:7 sicut fortes current quasi uiri bel-

wall like men of war; and they shall march every one on his ways, and they shall not break their ranks:

latores ascendent murum uir in uiis suis gradietur et non

2:8 Neither shall one thrust another; they shall walk every declinabunt a semitis suis 2:8 unusquisque fratrem suum one in his path: and when they fall upon the sword, they shall not be wounded.

non coartabit singuli in calle suo ambulabunt sed et per fe-

2:9 They shall run to and fro in the city; they shall run nestras cadent et non demolientur 2:9 urbem ingredientur

upon the wall, they shall climb up upon the houses; they shall enter in at the windows like a thief.

in muro current domos conscendent per fenestras intrabunt

2:10 The earth shall quake before them; the heavens shall quasi fur 2:10 a facie eius contremuit terra moti sunt caeli

tremble: the sun and the moon shall be dark, and the stars shall withdraw their shining:

sol et luna obtenebrati sunt et stellae retraxerunt splendorem

2:11 And the LORD shall utter his voice before his army: suum 2:11 et Dominus dedit uocem suam ante faciem exfor his camp is very great: for he is strong that executeth his word: for the day of the LORD is great and very terri-

ble; and who can abide it? ercitus sui quia multa sunt nimis castra eius quia fortia et

facientia uerbum eius magnus enim dies Domini et terribilis

2:12 Therefore also now, saith the LORD, turn ye even to ualde et quis sustinebit eum 2:12 nunc ergo dicit Dominus

me with all your heart, and with fasting, and with weeping, and with mourning:

conuertimini ad me in toto corde uestro in ieiunio et in fletu

2:13 And rend your heart, and not your garments, and turn et in planctu 2:13 et scindite corda uestra et non uestimenta

unto the LORD your God: for he is gracious and merciful, slow to anger, and of great kindness, and repenteth him of

the evil. uestra et conuertimini ad Dominum Deum uestrum quia be-

Prophetia Ioel

nignus et misericors est patiens et multae misericordiae et praestabilis super malitia 2:14 quis scit si conuertatur et ig- 2:14 Who knoweth if he will return and repent, and leave

a blessing behind him; even a meat offering and a drink offering unto the LORD your God?

noscat et relinquat post se benedictionem sacrificium et libamen Domino Deo nostro 2:15 canite tuba in Sion sanc- 2:15 Blow the trumpet in Zion, sanctify a fast, call a

solemn assembly:

tificate ieiunium uocate coetum 2:16 congregate populum 2:16 Gather the people, sanctify the congregation, assemble the elders, gather the children, and those that suck the breasts: let the bridegroom go forth of his chamber, and

sanctificate ecclesiam coadunate senes congregate paruulos the bride out of her closet. et sugentes ubera egrediatur sponsus de cubili suo et sponsa de thalamo suo 2:17 inter uestibulum et altare plorabunt 2:17 Let the priests, the ministers of the LORD, weep be-

tween the porch and the altar, and let them say, Spare thy people, O LORD, and give not thine heritage to reproach, that the heathen should rule over them: wherefore should they say among the people, Where is their God?

sacerdotes ministri Domini et dicent parce Domine populo tuo et ne des hereditatem tuam in obprobrium ut dominentur eis nationes quare dicunt in populis ubi est Deus eorum 2:18 zelatus est Dominus terram suam et pepercit populo suo 2:18 Then will the LORD be jealous for his land, and pity

his people.

2:19 et respondit Dominus et dixit populo suo ecce ego mit- 2:19 Yea, the LORD will answer and say unto his people,

Behold, I will send you corn, and wine, and oil, and ye shall be satisfied therewith: and I will no more make you

tam uobis frumentum et uinum et oleum et replebimini eo a reproach among the heathen: et non dabo uos ultra obprobrium in gentibus 2:20 et eum 2:20 But I will remove far off from you the northern army,

and will drive him into a land barren and desolate, with his face toward the east sea, and his hinder part toward the utmost sea, and his stink shall come up, and his ill savour shall come up, because he hath done great things.

qui ab aquilone est procul faciam a uobis et expellam eum in terram inuiam et desertam faciem eius contra mare orientale et extremum eius ad mare nouissimum et ascendet fetor eius et ascendet putredo eius quia superbe egit 2:21 noli timere 2:21 Fear not, O land; be glad and rejoice: for the LORD

will do great things.

terra exulta et laetare quoniam magnificauit Dominus ut faceret 2:22 nolite timere animalia regionis quia germinaue- 2:22 Be not afraid, ye beasts of the field: for the pastures

of the wilderness do spring, for the tree beareth her fruit, the fig tree and the vine do yield their strength.

runt speciosa deserti quia lignum adtulit fructum suum ficus et uinea dederunt uirtutem suam 2:23 et filii Sion exultate et 2:23 Be glad then, ye children of Zion, and rejoice in the

LORD your God: for he hath given you the former rain moderately, and he will cause to come down for you the

laetamini in Domino Deo uestro quia dedit uobis doctorem rain, the former rain, and the latter rain in the first month.

iustitiae et descendere faciet ad uos imbrem matutinum et serotinum in principio 2:24 et implebuntur areae frumento 2:24 And the floors shall be full of wheat, and the vats

shall overflow with wine and oil.

et redundabunt torcularia uino et oleo 2:25 et reddam uo- 2:25 And I will restore to you the years that the locust

hath eaten, the cankerworm, and the caterpiller, and the palmerworm, my great army which I sent among you.

bis annos quos comedit lucusta bruchus et rubigo et eruca

fortitudo mea magna quam misi in uos 2:26 et comedetis 2:26 And ye shall eat in plenty, and be satisfied, and praise the name of the LORD your God, that hath dealt wondrously with you: and my people shall never be ashamed.

uescentes et saturabimini et laudabitis nomen Domini Dei uestri qui fecit uobiscum mirabilia et non confundetur populus meus in sempiternum 2:27 et scietis quia in medio Isra- 2:27 And ye shall know that I am in the midst of Israel,

and that I am the LORD your God, and none else: and my people shall never be ashamed.

hel ego sum et ego Dominus Deus uester et non est amplius

Biblia Sacra Vulgata

2:28 And it shall come to pass afterward, that I will pour et non confundetur populus meus in aeternum 2:28 et erit out my spirit upon all flesh; and your sons and your daughters shall prophesy, your old men shall dream dreams, your

young men shall see visions: post haec effundam spiritum meum super omnem carnem

et prophetabunt filii uestri et filiae uestrae senes uestri som-

2:29 And also upon the servants and upon the handmaids nia somniabunt et iuuenes uestri uisiones uidebunt 2:29 sed

in those days will I pour out my spirit.

et super seruos et ancillas in diebus illis effundam spiritum

2:30 And I will shew wonders in the heavens and in the meum 2:30 et dabo prodiiia in caelo et in terra sanguinem

earth, blood, and fire, and pillars of smoke.

2:31 The sun shall be turned into darkness, and the moon et ignem et uaporem fumi 2:31 sol uertetur in tenebras et

into blood, before the great and terrible day of the LORD come.

luna in sanguinem antequam ueniat dies Domini magnus et

2:32 And it shall come to pass, that whosoever shall call horribilis 2:32 et erit omnis qui inuocauerit nomen Domini

on the name of the LORD shall be delivered: for in mount Zion and in Jerusalem shall be deliverance, as the LORD

hath said, and in the remnant whom the LORD shall call. saluus erit quia in monte Sion et in Hierusalem erit saluatio sicut dixit Dominus et in residuis quos Dominus uocauerit

3:1 For, behold, in those days, and in that time, when I 3:1 quia ecce in diebus illis et in tempore illo cum conuer-

shall bring again the captivity of Judah and Jerusalem, 3:2 I will also gather all nations, and will bring them down tero captiuitatem Iuda et Hierusalem 3:2 congregabo ominto the valley of Jehoshaphat, and will plead with them there for my people and for my heritage Israel, whom they

have scattered among the nations, and parted my land. nes gentes et deducam eas in ualle Iosaphat et disceptabo

cum eis ibi super populo meo et hereditate mea Israhel quos

3:3 And they have cast lots for my people; and have given disperserunt in nationibus et terram meam diuiserunt 3:3 et

a boy for an harlot, and sold a girl for wine, that they might drink.

super populum meum miserunt sortem et posuerunt puerum in prostibulum et puellam uendiderunt pro uino ut biberent

3:4 Yea, and what have ye to do with me, O Tyre, and Zi- 3:4 uerum quid uobis et mihi Tyrus et Sidon et omnis termi-

don, and all the coasts of Palestine? will ye render me a recompence? and if ye recompense me, swiftly and speed-

ily will I return your recompence upon your own head; nus Palestinorum numquid ultionem uos redditis mihi et si

ulciscimini uos contra me cito uelociter reddam uicissitudi-

3:5 Because ye have taken my silver and my gold, and nem uobis super caput uestrum 3:5 argentum enim meum

have carried into your temples my goodly pleasant things:

et aurum tulistis et desiderabilia mea et pulcherrima intuli-

3:6 The children also of Judah and the children of stis in delubra uestra 3:6 et filios Iuda et filios Hierusalem

Jerusalem have ye sold unto the Grecians, that ye might remove them far from their border.

uendidistis filiis Graecorum ut longe faceretis eos de finibus

3:7 Behold, I will raise them out of the place whither ye suis 3:7 ecce ego suscitabo eos de loco in quo uendidishave sold them, and will return your recompence upon your own head:

tis eos et conuertam retributionem uestram in caput uestrum

3:8 And I will sell your sons and your daughters into the 3:8 et uendam filios uestros et filias uestras in manibus filio-

hand of the children of Judah, and they shall sell them to the Sabeans, to a people far off: for the LORD hath spoken

it. rum Iuda et uenundabunt eos Sabeis genti longinquae quia 3:9 Proclaim ye this among the Gentiles; Prepare war, Dominus locutus est 3:9 clamate hoc in gentibus sancti-

wake up the mighty men, let all the men of war draw near; let them come up:

ficate bellum suscitate robustos accedant ascendant omnes

3:10 Beat your plowshares into swords and your prun- uiri bellatores 3:10 concidite aratra uestra in gladios et li-

inghooks into spears: let the weak say, I am strong.

gones uestros in lanceas infirmus dicat quia fortis ego sum

Prophetia Ioel

3:11 erumpite et uenite omnes gentes de circuitu et congrega- 3:11 Assemble yourselves, and come, all ye heathen, and

gather yourselves together round about: thither cause thy mighty ones to come down, O LORD.

mini ibi occumbere faciet Dominus robustos tuos 3:12 con-

3:12 Let the heathen be wakened, and come up to the valley of Jehoshaphat: for there will I sit to judge all the heathen round about.

surgant et ascendant gentes in uallem Iosaphat quia ibi sedebo ut iudicem omnes gentes in circuitu 3:13 mittite falces 3:13 Put ye in the sickle, for the harvest is ripe: come, get

you down; for the press is full, the fats overflow; for their wickedness is great.

quoniam maturauit messis uenite et descendite quia plenum est torcular exuberant torcularia quia multiplicata est malitia eorum 3:14 populi populi in ualle concisionis quia iuxta 3:14 Multitudes, multitudes in the valley of decision: for

the day of the LORD is near in the valley of decision.

est dies Domini in ualle concisionis 3:15 sol et luna obte- 3:15 The sun and the moon shall be darkened, and the stars

shall withdraw their shining.

nebricata sunt et stellae retraxerunt splendorem suum 3:16 3:16 The LORD also shall roar out of Zion, and utter his

voice from Jerusalem; and the heavens and the earth shall shake: but the LORD will be the hope of his people, and

et Dominus de Sion rugiet et de Hierusalem dabit uocem the strength of the children of Israel. suam et mouebuntur caeli et terra et Dominus spes populi sui et fortitudo filiorum Israhel 3:17 et scietis quia ego Do- 3:17 So shall ye know that I am the LORD your God

dwelling in Zion, my holy mountain: then shall Jerusalem be holy, and there shall no strangers pass through her any

minus Deus uester habitans in Sion in monte sancto meo et more. erit Hierusalem sancta et alieni non transibunt per eam amplius 3:18 et erit in die illa stillabunt montes dulcedinem 3:18 And it shall come to pass in that day, that the moun-

tains shall drop down new wine, and the hills shall flow with milk, and all the rivers of Judah shall flow with waters, and a fountain shall come forth out of the house of the LORD, and shall water the valley of Shittim.

et colles fluent lacte et per omnes riuos Iuda ibunt aquae et fons de domo Domini egredietur et inrigabit torrentem Spinarum 3:19 Aegyptus in desolatione erit et Idumea in de- 3:19 Egypt shall be a desolation, and Edom shall be a

desolate wilderness, for the violence against the children of Judah, because they have shed innocent blood in their

sertum perditionis pro eo quod inique egerint in filios Iuda land. et effuderint sanguinem innocentem in terra sua 3:20 et Iu- 3:20 But Judah shall dwell for ever, and Jerusalem from

generation to generation.

daea in aeternum habitabitur et Hierusalem in generatione

et generationem 3:21 et mundabo sanguinem eorum quem 3:21 For I will cleanse their blood that I have not cleansed:

for the LORD dwelleth in Zion.

non mundaueram et Dominus commorabitur in Sion

Biblia Sacra Vulgata

MCCLXXX

Prophetia Amos

1:1 uerba Amos qui fuit in pastoralibus de Thecuae quae 1:1 The words of Amos, who was among the herdmen

of Tekoa, which he saw concerning Israel in the days of Uzziah king of Judah, and in the days of Jeroboam the son

uidit super Israhel in diebus Oziae regis Iuda et in diebus of Joash king of Israel, two years before the earthquake. Hieroboam filii Ioas regis Israhel ante duos annos terraemotus 1:2 et dixit Dominus de Sion rugiet et de Hierusalem 1:2 And he said, The LORD will roar from Zion, and ut-

ter his voice from Jerusalem; and the habitations of the shepherds shall mourn, and the top of Carmel shall wither.

dabit uocem suam et luxerunt speciosa pastorum et exsiccatus est uertex Carmeli 1:3 haec dicit Dominus super tribus 1:3 Thus saith the LORD; For three transgressions of

Damascus, and for four, I will not turn away the punishment thereof; because they have threshed Gilead with

sceleribus Damasci et super quattuor non conuertam eum threshing instruments of iron: eo quod triturauerint in plaustris ferreis Galaad 1:4 et mit- 1:4 But I will send a fire into the house of Hazael, which

shall devour the palaces of Benhadad.

tam ignem in domum Azahel et deuorabit domos Benadad 1:5 et conteram uectem Damasci et disperdam habitatorem 1:5 I will break also the bar of Damascus, and cut off the

inhabitant from the plain of Aven, and him that holdeth the sceptre from the house of Eden: and the people of Syria

de campo Idoli et tenentem sceptrum de domo Voluptatis shall go into captivity unto Kir, saith the LORD. et transferetur populus Syriae Cyrenen dicit Dominus 1:6 1:6 Thus saith the LORD; For three transgressions of

Gaza, and for four, I will not turn away the punishment thereof; because they carried away captive the whole cap-

haec dicit Dominus super tribus sceleribus Gazae et super tivity, to deliver them up to Edom: quattuor non conuertam eum eo quod transtulerit captiuitatem perfectam ut concluderet eam in Idumea 1:7 et mit- 1:7 But I will send a fire on the wall of Gaza, which shall

devour the palaces thereof:

tam ignem in murum Gazae et deuorabit aedes eius 1:8 et 1:8 And I will cut off the inhabitant from Ashdod, and

him that holdeth the sceptre from Ashkelon, and I will turn mine hand against Ekron: and the remnant of the

disperdam habitatorem de Azoto et tenentem sceptrum de Philistines shall perish, saith the Lord GOD. Ascalone et conuertam manum meam super Accaron et peribunt reliqui Philisthinorum dicit Dominus Deus 1:9 haec 1:9 Thus saith the LORD; For three transgressions of

Tyrus, and for four, I will not turn away the punishment thereof; because they delivered up the whole captivity to

dicit Dominus super tribus sceleribus Tyri et super quattuor Edom, and remembered not the brotherly covenant: non conuertam eum eo quod concluserint captiuitatem perfectam in Idumea et non sint recordati foederis fratrum 1:10 1:10 But I will send a fire on the wall of Tyrus, which shall

devour the palaces thereof.

et emittam ignem in murum Tyri et deuorabit aedes eius 1:11 haec dicit Dominus super tribus sceleribus Edom et su- 1:11 Thus saith the LORD; For three transgressions of

Edom, and for four, I will not turn away the punishment thereof; because he did pursue his brother with the sword, and did cast off all pity, and his anger did tear perpetually, and he kept his wrath for ever:

per quattuor non conuertam eum eo quod persecutus sit in gladio fratrem suum et uiolauerit misericordiam eius et tenuerit ultra furorem suum et indignationem suam seruauerit

usque in finem 1:12 mittam ignem in Theman et deuorabit 1:12 But I will send a fire upon Teman, which shall devour

the palaces of Bozrah.

aedes Bosrae 1:13 haec dicit Dominus super tribus sceleri- 1:13 Thus saith the LORD; For three transgressions of the

children of Ammon, and for four, I will not turn away the punishment thereof; because they have ripped up the women with child of Gilead, that they might enlarge their border:

bus filiorum Ammon et super quattuor non conuertam eum eo quod dissecuerit praegnantes Galaad ad dilatandum ter-

minum suum 1:14 et succendam ignem in muro Rabbae et 1:14 But I will kindle a fire in the wall of Rabbah, and it shall devour the palaces thereof, with shouting in the day of battle, with a tempest in the day of the whirlwind:

deuorabit aedes eius in ululatu in die belli et in turbine in die

Biblia Sacra Vulgata

1:15 And their king shall go into captivity, he and his commotionis 1:15 et ibit Melchom in captiuitatem ipse et

princes together, saith the LORD.

principes eius simul dicit Dominus

2:1 Thus saith the LORD; For three transgressions of 2:1 haec dicit Dominus super tribus sceleribus Moab et

Moab, and for four, I will not turn away the punishment thereof; because he burned the bones of the king of Edom

into lime: super quattuor non conuertam eum eo quod incenderit ossa

2:2 But I will send a fire upon Moab, and it shall devour regis Idumeae usque ad cinerem 2:2 et mittam ignem in the palaces of Kirioth: and Moab shall die with tumult, with shouting, and with the sound of the trumpet:

Moab et deuorabit aedes Carioth et morietur in sonitu Moab

2:3 And I will cut off the judge from the midst thereof, and in clangore tubae 2:3 et disperdam iudicem de medio eius will slay all the princes thereof with him, saith the LORD.

et omnes principes eius interficiam cum eo dicit Dominus

2:4 Thus saith the LORD; For three transgressions of 2:4 haec dicit Dominus super tribus sceleribus Iuda et su-

Judah, and for four, I will not turn away the punishment thereof; because they have despised the law of the LORD, and have not kept his commandments, and their lies caused them to err, after the which their fathers have walked:

per quattuor non conuertam eum eo quod abiecerint legem Domini et mandata eius non custodierint deceperunt enim

2:5 But I will send a fire upon Judah, and it shall devour eos idola sua post quae abierant patres eorum 2:5 et mit-

the palaces of Jerusalem. 2:6 Thus saith the LORD; For three transgressions of Is- tam ignem in Iuda et deuorabit aedes Hierusalem 2:6 haec rael, and for four, I will not turn away the punishment thereof; because they sold the righteous for silver, and the

poor for a pair of shoes; dicit Dominus super tribus sceleribus Israhel et super quat-

tuor non conuertam eum pro eo quod uendiderit argento iu-

2:7 That pant after the dust of the earth on the head of the stum et pauperem pro calciamentis 2:7 qui conterunt super

poor, and turn aside the way of the meek: and a man and his father will go in unto the same maid, to profane my

holy name: puluerem terrae capita pauperum et uiam humilium decli-

nant et filius ac pater eius ierunt ad puellam ut uiolarent no-

2:8 And they lay themselves down upon clothes laid to men sanctum meum 2:8 et super uestimentis pigneratis ac-

pledge by every altar, and they drink the wine of the condemned in the house of their god.

cubuerunt iuxta omne altare et uinum damnatorum bibebant

2:9 Yet destroyed I the Amorite before them, whose height in domo Dei sui 2:9 ego autem exterminaui Amorreum a was like the height of the cedars, and he was strong as the oaks; yet I destroyed his fruit from above, and his roots

from beneath. facie eorum cuius altitudo cedrorum altitudo eius et fortis

ipse quasi quercus et contriui fructum eius desuper et radi-

2:10 Also I brought you up from the land of Egypt, and ces eius subter 2:10 ego sum qui ascendere uos feci de terra

led you forty years through the wilderness, to possess the land of the Amorite.

Aegypti et eduxi uos in deserto quadraginta annis ut possi-

2:11 And I raised up of your sons for prophets, and of your deretis terram Amorrei 2:11 et suscitaui de filiis uestris in

young men for Nazarites. Is it not even thus, O ye children of Israel? saith the LORD.

prophetas et de iuuenibus uestris nazarenos numquid non ita

2:12 But ye gave the Nazarites wine to drink; and com- est filii Israhel dicit Dominus 2:12 et propinabatis nazarenis

manded the prophets, saying, Prophesy not.

2:13 Behold, I am pressed under you, as a cart is pressed uino et prophetis mandabatis dicentes ne prophetetis 2:13

that is full of sheaves.

ecce ego stridebo super uos sicut stridet plaustrum onustum

2:14 Therefore the flight shall perish from the swift, and faeno 2:14 et peribit fuga a ueloce et fortis non obtinebit

the strong shall not strengthen his force, neither shall the mighty deliver himself:

uirtutem suam et robustus non saluabit animam suam 2:15

2:15 Neither shall he stand that handleth the bow; and he that is swift of foot shall not deliver himself: neither shall he that rideth the horse deliver himself.

et tenens arcum non stabit et uelox pedibus suis non salua-

Prophetia Amos

bitur et ascensor equi non saluabit animam suam 2:16 et 2:16 And he that is courageous among the mighty shall

flee away naked in that day, saith the LORD.

robustus corde inter fortes nudus fugiet in die illa dicit Dominus

3:1 audite uerbum quod locutus est Dominus super uos filii 3:1 Hear this word that the LORD hath spoken against

you, O children of Israel, against the whole family which I brought up from the land of Egypt, saying,

Israhel super omni cognatione quam eduxi de terra Aegypti dicens 3:2 tantummodo uos cognoui ex omnibus cognatio- 3:2 You only have I known of all the families of the earth:

therefore I will punish you for all your iniquities.

nibus terrae idcirco uisitabo super uos omnes iniquitates uestras 3:3 numquid ambulabunt duo pariter nisi conuenerit 3:3 Can two walk together, except they be agreed? eis 3:4 numquid rugiet leo in saltu nisi habuerit praedam 3:4 Will a lion roar in the forest, when he hath no prey?

will a young lion cry out of his den, if he have taken nothing?

numquid dabit catulus leonis uocem de cubili suo nisi aliquid adprehenderit 3:5 numquid cadet auis in laqueum ter- 3:5 Can a bird fall in a snare upon the earth, where no gin

is for him? shall one take up a snare from the earth, and have taken nothing at all?

rae absque aucupe numquid auferetur laqueus de terra antequam quid ceperit 3:6 si clanget tuba in ciuitate et popu- 3:6 Shall a trumpet be blown in the city, and the people

not be afraid? shall there be evil in a city, and the LORD hath not done it?

lus non expauescet si erit malum in ciuitate quod Dominus non fecit 3:7 quia non faciet Dominus Deus uerbum nisi re- 3:7 Surely the Lord GOD will do nothing, but he revealeth

his secret unto his servants the prophets.

uelauerit secretum suum ad seruos suos prophetas 3:8 leo 3:8 The lion hath roared, who will not fear? the Lord GOD

hath spoken, who can but prophesy?

rugiet quis non timebit Dominus Deus locutus est quis non prophetabit 3:9 auditum facite in aedibus Azoti et in aedi- 3:9 Publish in the palaces at Ashdod, and in the palaces in

the land of Egypt, and say, Assemble yourselves upon the mountains of Samaria, and behold the great tumults in the

bus terrae Aegypti et dicite congregamini super montes Sa- midst thereof, and the oppressed in the midst thereof. mariae et uidete insanias multas in medio eius et calumniam patientes in penetrabilibus eius 3:10 et nescierunt facere re- 3:10 For they know not to do right, saith the LORD, who

store up violence and robbery in their palaces.

ctum dicit Dominus thesaurizantes iniquitatem et rapinas in aedibus suis 3:11 propterea haec dicit Dominus Deus tribu- 3:11 Therefore thus saith the Lord GOD; An adversary

there shall be even round about the land; and he shall bring down thy strength from thee, and thy palaces shall

labitur et circumietur terra et detrahetur ex te fortitudo tua et be spoiled. diripientur aedes tuae 3:12 haec dicit Dominus quomodo si 3:12 Thus saith the LORD; As the shepherd taketh out of

the mouth of the lion two legs, or a piece of an ear; so shall the children of Israel be taken out that dwell in Samaria in

eruat pastor de ore leonis duo crura aut extremum auriculae the corner of a bed, and in Damascus in a couch. sic eruentur filii Israhel qui habitant in Samaria in plaga lectuli et in Damasco grabatti 3:13 audite et contestamini in 3:13 Hear ye, and testify in the house of Jacob, saith the

Lord GOD, the God of hosts,

domo Iacob dicit Dominus Deus exercituum 3:14 quia in 3:14 That in the day that I shall visit the transgressions of Israel upon him I will also visit the altars of Bethel: and the horns of the altar shall be cut off, and fall to the ground.

die cum uisitare coepero praeuaricationes Israhel super eum uisitabo et super altaria Bethel et amputabuntur cornua altaris et cadent in terram 3:15 et percutiam domum hiemalem 3:15 And I will smite the winter house with the summer

house; and the houses of ivory shall perish, and the great houses shall have an end, saith the LORD.

cum domo aestiua et peribunt domus eburneae et dissipabun-

Biblia Sacra Vulgata

tur aedes multae dicit Dominus

4:1 Hear this word, ye kine of Bashan, that are in the 4:1 audite uerbum hoc uaccae pingues quae estis in monte

mountain of Samaria, which oppress the poor, which crush the needy, which say to their masters, Bring, and let us

drink. Samariae quae calumniam facitis egenis et confringitis pau-4:2 The Lord GOD hath sworn by his holiness, that, lo, peres quae dicitis dominis uestris adferte et bibemus 4:2

the days shall come upon you, that he will take you away with hooks, and your posterity with fishhooks.

iurauit Dominus Deus in sancto suo quia ecce dies uenient super uos et leuabunt uos in contis et reliquias uestras in

4:3 And ye shall go out at the breaches, every cow at that ollis feruentibus 4:3 et per aperturas exibitis altera contra

which is before her; and ye shall cast them into the palace, saith the LORD.

alteram et proiciemini in Armon dicit Dominus 4:4 uenite

4:4 Come to Bethel, and transgress; at Gilgal multiply transgression; and bring your sacrifices every morning, and your tithes after three years:

ad Bethel et impie agite ad Galgalam et multiplicate praeuaricationem et offerte mane uictimas uestras tribus diebus

4:5 And offer a sacrifice of thanksgiving with leaven, and decimas uestras 4:5 et sacrificate de fermentato laudem et

proclaim and publish the free offerings: for this liketh you, O ye children of Israel, saith the Lord GOD.

uocate uoluntarias oblationes et adnuntiate sic enim uoluistis

4:6 And I also have given you cleanness of teeth in all your filii Israhel dicit Dominus Deus 4:6 unde et ego dedi uobis cities, and want of bread in all your places: yet have ye not returned unto me, saith the LORD.

stuporem dentium in cunctis urbibus uestris et indigentiam panum in omnibus locis uestris et non estis reuersi ad me di-

4:7 And also I have withholden the rain from you, when cit Dominus 4:7 ego quoque prohibui a uobis imbrem cum

there were yet three months to the harvest: and I caused it to rain upon one city, and caused it not to rain upon another city: one piece was rained upon, and the piece whereupon it rained not withered.

adhuc tres menses superessent usque ad messem et plui super ciuitatem unam et super ciuitatem alteram non plui pars

4:8 So two or three cities wandered unto one city, to drink una conpluta est et pars super quam non plui aruit 4:8 et water; but they were not satisfied: yet have ye not returned unto me, saith the LORD.

uenerunt duae et tres ciuitates ad ciuitatem unam ut biberent aquam et non sunt satiatae et non rediitis ad me dicit Domi-

4:9 I have smitten you with blasting and mildew: when nus 4:9 percussi uos in uento urente et in aurugine multi-

your gardens and your vineyards and your fig trees and your olive trees increased, the palmerworm devoured

them: yet have ye not returned unto me, saith the LORD. tudinem hortorum uestrorum et uinearum uestrarum oliueta

uestra et ficeta uestra comedit eruca et non rediitis ad me

4:10 I have sent among you the pestilence after the manner dicit Dominus 4:10 misi in uos mortem in uia Aegypti perof Egypt: your young men have I slain with the sword, and have taken away your horses; and I have made the stink of your camps to come up unto your nostrils: yet have ye not returned unto me, saith the LORD.

cussi in gladio iuuenes uestros usque ad captiuitatem equorum uestrorum et ascendere feci putredinem castrorum uestrorum in nares uestras et non rediitis ad me dicit Dominus

4:11 I have overthrown some of you, as God overthrew 4:11 subuerti uos sicut subuertit Deus Sodomam et Gomor-

Sodom and Gomorrah, and ye were as a firebrand plucked out of the burning: yet have ye not returned unto me, saith

the LORD. ram et facti estis quasi torris raptus de incendio et non red-4:12 Therefore thus will I do unto thee, O Israel: and be- iitis ad me dicit Dominus 4:12 quapropter haec faciam tibi

cause I will do this unto thee, prepare to meet thy God, O Israel.

Israhel postquam autem haec fecero tibi praeparare in oc-

4:13 For, lo, he that formeth the mountains, and createth cursum Dei tui Israhel 4:13 quia ecce formans montes et the wind, and declareth unto man what is his thought, that maketh the morning darkness, and treadeth upon the high places of the earth, The LORD, The God of hosts, is his name.

Prophetia Amos

creans uentum et adnuntians homini eloquium suum faciens matutinam nebulam et gradiens super excelsa terrae Dominus Deus exercituum nomen eius

5:1 audite uerbum istud quod ego leuo super uos planc- 5:1 Hear ye this word which I take up against you, even a

lamentation, O house of Israel.

tum domus Israhel cecidit non adiciet ut resurgat 5:2 uirgo 5:2 The virgin of Israel is fallen; she shall no more rise:

she is forsaken upon her land; there is none to raise her up.

Israhel proiecta est in terram suam non est qui suscitet eam 5:3 quia haec dicit Dominus Deus urbs de qua egredieban- 5:3 For thus saith the Lord GOD; The city that went out

by a thousand shall leave an hundred, and that which went forth by an hundred shall leave ten, to the house of Israel.

tur mille relinquentur in ea centum et de qua egredieban-

tur centum relinquentur in ea decem in domo Israhel 5:4 5:4 For thus saith the LORD unto the house of Israel, Seek

ye me, and ye shall live:

quia haec dicit Dominus domui Israhel quaerite me et uiuetis 5:5 et nolite quaerere Bethel et in Galgala nolite in- 5:5 But seek not Bethel, nor enter into Gilgal, and pass not

to Beersheba: for Gilgal shall surely go into captivity, and Bethel shall come to nought.

trare et in Bersabee non transibitis quia Galgala captiua ducetur et Bethel erit inutilis 5:6 quaerite Dominum et uiuite 5:6 Seek the LORD, and ye shall live; lest he break out

like fire in the house of Joseph, and devour it, and there be none to quench it in Bethel.

ne forte conburatur ut ignis domus Ioseph et deuorabit et non erit qui extinguat Bethel 5:7 qui conuertitis in absin- 5:7 Ye who turn judgment to wormwood, and leave off

righteousness in the earth,

thium iudicium et iustitiam in terra relinquitis 5:8 facien- 5:8 Seek him that maketh the seven stars and Orion, and

turneth the shadow of death into the morning, and maketh the day dark with night: that calleth for the waters of the sea, and poureth them out upon the face of the earth: The LORD is his name:

tem Arcturum et Orionem et conuertentem in mane tenebras et diem nocte mutantem qui uocat aquas maris et effun-

dit eas super faciem terrae Dominus nomen eius 5:9 qui 5:9 That strengtheneth the spoiled against the strong, so

that the spoiled shall come against the fortress.

subridet uastitatem super robustum et depopulationem super potentem adfert 5:10 odio habuerunt in porta corripien- 5:10 They hate him that rebuketh in the gate, and they ab-

hor him that speaketh uprightly.

tem et loquentem perfecte abominati sunt 5:11 idcirco pro 5:11 Forasmuch therefore as your treading is upon the

poor, and ye take from him burdens of wheat: ye have built houses of hewn stone, but ye shall not dwell in them; ye have planted pleasant vineyards, but ye shall not drink wine of them.

eo quod diripiebatis pauperem et praedam electam tollebatis ab eo domos quadro lapide aedificabitis et non habitabitis in eis uineas amantissimas plantabitis et non bibetis uinum earum 5:12 quia cognoui multa scelera uestra et fortia 5:12 For I know your manifold transgressions and your

mighty sins: they afflict the just, they take a bribe, and they turn aside the poor in the gate from their right.

peccata uestra hostes iusti accipientes munus et pauperes in porta deprimentes 5:13 ideo prudens in tempore illo tacebit 5:13 Therefore the prudent shall keep silence in that time;

for it is an evil time.

quia tempus malum est 5:14 quaerite bonum et non malum 5:14 Seek good, and not evil, that ye may live: and so the LORD, the God of hosts, shall be with you, as ye have spoken.

ut uiuatis et erit Dominus Deus exercituum uobiscum sicut dixistis 5:15 odite malum et diligite bonum et constituite 5:15 Hate the evil, and love the good, and establish judg-

ment in the gate: it may be that the LORD God of hosts will be gracious unto the remnant of Joseph.

in porta iudicium si forte misereatur Dominus Deus exercituum reliquiis Ioseph 5:16 propterea haec dicit Dominus 5:16 Therefore the LORD, the God of hosts, the LORD,

saith thus; Wailing shall be in all streets; and they shall say in all the highways, Alas! alas! and they shall call the husbandman to mourning, and such as are skilful of lamentation to wailing.

Biblia Sacra Vulgata

Deus exercituum Dominator in omnibus plateis planctus et in cunctis quae foris sunt dicetur uae uae et uocabunt agrico-

5:17 And in all vineyards shall be wailing: for I will pass lam ad luctum et ad planctum eos qui sciunt plangere 5:17 through thee, saith the LORD.

et in omnibus uineis erit planctus quia pertransibo in medio

5:18 Woe unto you that desire the day of the LORD! to tui dicit Dominus 5:18 uae desiderantibus diem Domini ad

what end is it for you? the day of the LORD is darkness, and not light.

quid eam uobis dies Domini ista tenebrae et non lux 5:19

5:19 As if a man did flee from a lion, and a bear met him; or went into the house, and leaned his hand on the wall, and a serpent bit him.

quomodo si fugiat uir a facie leonis et occurrat ei ursus et ingrediatur domum et innitatur manu sua super parietem et

5:20 Shall not the day of the LORD be darkness, and not mordeat eum coluber 5:20 numquid non tenebrae dies Do-

light? even very dark, and no brightness in it? 5:21 I hate, I despise your feast days, and I will not smell mini et non lux et caligo et non splendor in ea 5:21 odi et in your solemn assemblies.

proieci festiuitates uestras et non capiam odorem coetuum

5:22 Though ye offer me burnt offerings and your meat uestrorum 5:22 quod si adtuleritis mihi holocaustomata et

offerings, I will not accept them: neither will I regard the peace offerings of your fat beasts.

munera uestra non suscipiam et uota pinguium uestrorum

5:23 Take thou away from me the noise of thy songs; for I non respiciam 5:23 aufer a me tumultum carminum tuorum

will not hear the melody of thy viols.

5:24 But let judgment run down as waters, and righteous- et cantica lyrae tuae non audiam 5:24 et reuelabitur quasi

ness as a mighty stream.

5:25 Have ye offered unto me sacrifices and offerings in aqua iudicium et iustitia quasi torrens fortis 5:25 numquid

the wilderness forty years, O house of Israel?

hostias et sacrificium obtulistis mihi in deserto quadraginta

5:26 But ye have borne the tabernacle of your Moloch and annis domus Israhel 5:26 et portastis tabernaculum Moloch

Chiun your images, the star of your god, which ye made to yourselves.

uestro et imaginem idolorum uestrorum sidus dei uestri quae

5:27 Therefore will I cause you to go into captivity beyond fecistis uobis 5:27 et migrare uos faciam trans Damascum Damascus, saith the LORD, whose name is The God of hosts.

dixit Dominus Deus exercituum nomen eius

6:1 Woe to them that are at ease in Zion, and trust in the 6:1 uae qui opulenti estis in Sion et confiditis in monte Sa-

mountain of Samaria, which are named chief of the nations, to whom the house of Israel came!

mariae optimates capita populorum ingredientes pompatice

6:2 Pass ye unto Calneh, and see; and from thence go ye to domum Israhel 6:2 transite in Chalanne et uidete et ite inde

Hamath the great: then go down to Gath of the Philistines: be they better than these kingdoms? or their border greater

than your border? in Emath magnam et descendite in Geth Palestinorum et ad

optima quaeque regna horum si latior terminus eorum ter-

6:3 Ye that put far away the evil day, and cause the seat of mino uestro est 6:3 qui separati estis in diem malum et ad-

violence to come near;

6:4 That lie upon beds of ivory, and stretch themselves propinquatis solio iniquitatis 6:4 qui dormitis in lectis ebur-

upon their couches, and eat the lambs out of the flock, and the calves out of the midst of the stall;

neis et lasciuitis in stratis uestris qui comeditis agnum de

6:5 That chant to the sound of the viol, and invent to them- grege et uitulos de medio armenti 6:5 qui canitis ad uocem selves instruments of musick, like David;

6:6 That drink wine in bowls, and anoint themselves with psalterii sicut Dauid putauerunt se habere uasa cantici 6:6

the chief ointments: but they are not grieved for the affliction of Joseph.

bibentes in fialis uinum et optimo unguento delibuti et nihil

6:7 Therefore now shall they go captive with the first that patiebantur super contritione Ioseph 6:7 quapropter nunc

go captive, and the banquet of them that stretched themselves shall be removed.

Prophetia Amos

migrabunt in capite transmigrantium et auferetur factio lasciuientium 6:8 iurauit Dominus Deus in anima sua dicit 6:8 The Lord GOD hath sworn by himself, saith the LORD

the God of hosts, I abhor the excellency of Jacob, and hate his palaces: therefore will I deliver up the city with all that

Dominus Deus exercituum detestor ego superbiam Iacob et is therein. domos eius odi et tradam ciuitatem cum habitatoribus suis 6:9 quod si reliqui fuerint decem uiri in domo una et ipsi mo- 6:9 And it shall come to pass, if there remain ten men in

one house, that they shall die.

rientur 6:10 et tollet eum propinquus suus et conburet eum 6:10 And a man’s uncle shall take him up, and he that

burneth him, to bring out the bones out of the house, and shall say unto him that is by the sides of the house, Is there

ut efferat ossa de domo et dicet ei qui in penetrabilibus do- yet any with thee? mus est numquid adhuc est apud te 6:11 et respondebit finis 6:11 and he shall say, No. Then shall he say, Hold thy

tongue: for we may not make mention of the name of the LORD.

est et dicet ei tace et non recorderis nominis Domini 6:12

6:12 For, behold, the LORD commandeth, and he will smite the great house with breaches, and the little house with clefts.

quia ecce Dominus mandabit et percutiet domum maiorem ruinis et domum minorem scissionibus 6:13 numquid cur- 6:13 Shall horses run upon the rock? will one plow there

with oxen? for ye have turned judgment into gall, and the fruit of righteousness into hemlock:

rere queunt in petris equi aut arari potest in bubalis quoniam conuertistis in amaritudinem iudicium et fructum iustitiae in absinthium 6:14 qui laetamini in nihili qui dicitis numquid 6:14 Ye which rejoice in a thing of nought, which say,

Have we not taken to us horns by our own strength?

non in fortitudine nostra adsumpsimus nobis cornua 6:15 6:15 But, behold, I will raise up against you a nation, O

house of Israel, saith the LORD the God of hosts; and they shall afflict you from the entering in of Hemath unto the

ecce enim suscitabo super uos domus Israhel dicit Dominus river of the wilderness. Deus exercituum gentem et conterent uos ab introitu Emath usque ad torrentem Deserti

7:1 haec ostendit mihi Dominus Deus et ecce fictor lucus- 7:1 Thus hath the Lord GOD shewed unto me; and, be-

hold, he formed grasshoppers in the beginning of the shooting up of the latter growth; and, lo, it was the latter

tae in principio germinantium serotini imbris et ecce seroti- growth after the king’s mowings. nus post tonsorem regis 7:2 et factum est cum consummas- 7:2 And it came to pass, that when they had made an end

of eating the grass of the land, then I said, O Lord GOD, forgive, I beseech thee: by whom shall Jacob arise? for he

set comedere herbam terrae et dixi Domine Deus propitius is small. esto obsecro quis suscitabit Iacob quia paruulus est 7:3 mi- 7:3 The LORD repented for this: It shall not be, saith the

LORD.

sertus est Dominus super hoc non erit dixit Dominus 7:4 7:4 Thus hath the Lord GOD shewed unto me: and, be-

hold, the Lord GOD called to contend by fire, and it devoured the great deep, and did eat up a part.

haec ostendit mihi Dominus Deus et ecce uocabat iudicium ad ignem Dominus Deus et deuorauit abyssum multam et comedit simul partem 7:5 et dixi Domine Deus quiesce ob- 7:5 Then said I, O Lord GOD, cease, I beseech thee: by

whom shall Jacob arise? for he is small.

secro quis suscitabit Iacob quia paruulus est 7:6 misertus 7:6 The LORD repented for this: This also shall not be,

saith the Lord GOD.

est Dominus super hoc sed et istud non erit dixit Dominus Deus 7:7 haec ostendit mihi et ecce Dominus stans super 7:7 Thus he shewed me: and, behold, the LORD stood

upon a wall made by a plumbline, with a plumbline in his hand.

murum litum et in manu eius trulla cementarii 7:8 et dixit

7:8 And the LORD said unto me, Amos, what seest thou? And I said, A plumbline. Then said the LORD, Behold, I will set a plumbline in the midst of my people Israel: I will not again pass by them any more:

Dominus ad me quid tu uides Amos et dixi trullam cementarii et dixit Dominus ecce ego ponam trullam in medio populi

Biblia Sacra Vulgata

7:9 And the high places of Isaac shall be desolate, and mei Israhel non adiciam ultra superinducere eum 7:9 et de-

the sanctuaries of Israel shall be laid waste; and I will rise against the house of Jeroboam with the sword.

molientur excelsa idoli et sanctificationes Israhel desolabun-

7:10 Then Amaziah the priest of Bethel sent to Jeroboam tur et consurgam super domum Hieroboam in gladio 7:10 king of Israel, saying, Amos hath conspired against thee in the midst of the house of Israel: the land is not able to

bear all his words. et misit Amasias sacerdos Bethel ad Hieroboam regem Isra-

hel dicens rebellauit contra te Amos in medio domus Isra-

7:11 For thus Amos saith, Jeroboam shall die by the hel non poterit terra sustinere uniuersos sermones eius 7:11

sword, and Israel shall surely be led away captive out of their own land.

haec enim dicit Amos in gladio morietur Hieroboam et Isra-

7:12 Also Amaziah said unto Amos, O thou seer, go, flee hel captiuus migrabit de terra sua 7:12 et dixit Amasias ad

thee away into the land of Judah, and there eat bread, and prophesy there:

Amos qui uides gradere fuge in terram Iuda et comede ibi

7:13 But prophesy not again any more at Bethel: for it is panem et ibi prophetabis 7:13 et in Bethel non adicies ultra

the king’s chapel, and it is the king’s court.

ut prophetes quia sanctificatio regis est et domus regni est

7:14 Then answered Amos, and said to Amaziah, I was 7:14 et respondit Amos et dixit ad Amasiam non sum pro-

no prophet, neither was I a prophet’s son; but I was an herdman, and a gatherer of sycomore fruit:

pheta et non sum filius prophetae sed armentarius ego sum

7:15 And the LORD took me as I followed the flock, and uellicans sycomoros 7:15 et tulit me Dominus cum seque-

the LORD said unto me, Go, prophesy unto my people Israel.

rer gregem et dixit ad me Dominus uade propheta ad po-

7:16 Now therefore hear thou the word of the LORD: pulum meum Israhel 7:16 et nunc audi uerbum Domini tu

Thou sayest, Prophesy not against Israel, and drop not thy word against the house of Isaac.

dicis non prophetabis super Israhel et non stillabis super do-

7:17 Therefore thus saith the LORD; Thy wife shall be an mum idoli 7:17 propter hoc haec dicit Dominus uxor tua in

harlot in the city, and thy sons and thy daughters shall fall by the sword, and thy land shall be divided by line; and thou shalt die in a polluted land: and Israel shall surely go into captivity forth of his land.

ciuitate fornicabitur et filii tui et filiae tuae in gladio cadent et humus tua funiculo metietur et tu in terra polluta morieris et Israhel captiuus migrabit de terra sua

8:1 Thus hath the Lord GOD shewed unto me: and behold 8:1 haec ostendit mihi Dominus Deus et ecce uncinus po-

a basket of summer fruit.

8:2 And he said, Amos, what seest thou? And I said, A morum 8:2 et dixit quid tu uides Amos et dixi uncinum po-

basket of summer fruit. Then said the LORD unto me, The end is come upon my people of Israel; I will not again

pass by them any more. morum et dixit Dominus ad me uenit finis super populum

8:3 And the songs of the temple shall be howlings in that meum Israhel non adiciam ultra ut pertranseam eum 8:3 day, saith the Lord GOD: there shall be many dead bodies in every place; they shall cast them forth with silence.

et stridebunt cardines templi in die illa dicit Dominus Deus

8:4 Hear this, O ye that swallow up the needy, even to multi morientur in omni loco proicietur silentium 8:4 au-

make the poor of the land to fail,

dite hoc qui conteritis pauperem et deficere facitis egenos

8:5 Saying, When will the new moon be gone, that we terrae 8:5 dicentes quando transibit mensis et uenundabi-

may sell corn? and the sabbath, that we may set forth wheat, making the ephah small, and the shekel great, and

falsifying the balances by deceit? mus merces et sabbatum et aperiemus frumentum ut inmi-

nuamus mensuram et augeamus siclum et subponamus state-

8:6 That we may buy the poor for silver, and the needy for ras dolosas 8:6 ut possideamus in argento egenos et paupe-

a pair of shoes; yea, and sell the refuse of the wheat?

8:7 The LORD hath sworn by the excellency of Jacob, res pro calciamentis et quisquilias frumenti uendamus 8:7

Surely I will never forget any of their works.

Prophetia Amos

iurauit Dominus in superbia Iacob si oblitus fuero usque ad finem omnia opera eorum 8:8 numquid super isto non com- 8:8 Shall not the land tremble for this, and every one

mourn that dwelleth therein? and it shall rise up wholly as a flood; and it shall be cast out and drowned, as by the

mouebitur terra et lugebit omnis habitator eius et ascendet flood of Egypt. quasi fluuius uniuersus et eicietur et defluet quasi riuus Aegypti 8:9 et erit in die illa dicit Dominus occidet sol meridie 8:9 And it shall come to pass in that day, saith the Lord

GOD, that I will cause the sun to go down at noon, and I will darken the earth in the clear day:

et tenebrescere faciam terram in die luminis 8:10 et conuer-

8:10 And I will turn your feasts into mourning, and all your songs into lamentation; and I will bring up sackcloth upon all loins, and baldness upon every head; and I will make it as the mourning of an only son, and the end thereof as a bitter day.

tam festiuitates uestras in luctum et omnia cantica uestra in planctum et inducam super omne dorsum uestrum saccum et super omne caput caluitium et ponam eam quasi luctum unigeniti et nouissima eius quasi diem amarum 8:11 ecce dies 8:11 Behold, the days come, saith the Lord GOD, that I

will send a famine in the land, not a famine of bread, nor a thirst for water, but of hearing the words of the LORD:

ueniunt dicit Dominus et mittam famem in terram non famem panis neque sitim aquae sed audiendi uerbum Domini 8:12 et commouebuntur a mari usque ad mare et ab aquilone 8:12 And they shall wander from sea to sea, and from the

north even to the east, they shall run to and fro to seek the word of the LORD, and shall not find it.

usque ad orientem circumibunt quaerentes uerbum Domini et non inuenient 8:13 in die illa deficient uirgines pulchrae 8:13 In that day shall the fair virgins and young men faint

for thirst.

et adulescentes in siti 8:14 qui iurant in delicto Samariae et 8:14 They that swear by the sin of Samaria, and say, Thy

god, O Dan, liveth; and, The manner of Beersheba liveth; even they shall fall, and never rise up again.

dicunt uiuit deus tuus Dan et uiuit uia Bersabee et cadent et non resurgent ultra

9:1 uidi Dominum stantem super altare et dixit percute 9:1 I saw the LORD standing upon the altar: and he said,

Smite the lintel of the door, that the posts may shake: and cut them in the head, all of them; and I will slay the last of them with the sword: he that fleeth of them shall not flee away, and he that escapeth of them shall not be delivered.

cardinem et commoueantur superliminaria auaritia enim in capite omnium et nouissimum eorum in gladio interficiam non erit fuga eis fugiet et non saluabitur ex eis qui fuge-

rit 9:2 si descenderint usque ad infernum inde manus mea 9:2 Though they dig into hell, thence shall mine hand take them; though they climb up to heaven, thence will I bring them down:

educet eos et si ascenderint usque ad caelum inde detraham eos 9:3 et si absconditi fuerint in uertice Carmeli inde scru- 9:3 And though they hide themselves in the top of Carmel,

I will search and take them out thence; and though they be hid from my sight in the bottom of the sea, thence will I

tans auferam eos et si celauerint se ab oculis meis in fundo command the serpent, and he shall bite them: maris ibi mandabo serpenti et mordebit eos 9:4 et si abie- 9:4 And though they go into captivity before their ene-

mies, thence will I command the sword, and it shall slay them: and I will set mine eyes upon them for evil, and not

rint in captiuitatem coram inimicis suis ibi mandabo gladio for good. et occidet eos et ponam oculos meos super eos in malum et non in bonum 9:5 et Dominus Deus exercituum qui tangit 9:5 And the Lord GOD of hosts is he that toucheth the

land, and it shall melt, and all that dwell therein shall mourn: and it shall rise up wholly like a flood; and shall

terram et tabescet et lugebunt omnes habitantes in ea et as- be drowned, as by the flood of Egypt. cendet sicut riuus omnis et defluet sicut fluuius Aegypti 9:6 9:6 It is he that buildeth his stories in the heaven, and hath

founded his troop in the earth; he that calleth for the waters of the sea, and poureth them out upon the face of the earth:

qui aedificat in caelo ascensionem suam et fasciculum suum The LORD is his name.

Biblia Sacra Vulgata

super terram fundauit qui uocat aquas maris et effundit eas

9:7 Are ye not as children of the Ethiopians unto me, O super faciem terrae Dominus nomen eius 9:7 numquid non

children of Israel? saith the LORD. Have not I brought up Israel out of the land of Egypt? and the Philistines from

Caphtor, and the Syrians from Kir? ut filii Aethiopum uos estis mihi filii Israhel ait Dominus

numquid non Israhel ascendere feci de terra Aegypti et Pa-

9:8 Behold, the eyes of the Lord GOD are upon the sin- lestinos de Cappadocia et Syros de Cyrene 9:8 ecce oculi

ful kingdom, and I will destroy it from off the face of the earth; saving that I will not utterly destroy the house of

Jacob, saith the LORD. Domini Dei super regnum peccans et conteram illud a facie

terrae uerumtamen conterens non conteram domum Iacob

9:9 For, lo, I will command, and I will sift the house of dicit Dominus 9:9 ecce enim ego mandabo et concutiam in

Israel among all nations, like as corn is sifted in a sieve, yet shall not the least grain fall upon the earth.

omnibus gentibus domum Israhel sicut concutitur in cribro

9:10 All the sinners of my people shall die by the sword, et non cadet lapillus super terram 9:10 in gladio morientur

which say, The evil shall not overtake nor prevent us.

omnes peccatores populi mei qui dicunt non adpropinqua-

9:11 In that day will I raise up the tabernacle of David that bit et non ueniet super nos malum 9:11 in die illo suscitabo is fallen, and close up the breaches thereof; and I will raise up his ruins, and I will build it as in the days of old:

tabernaculum Dauid quod cecidit et reaedificabo aperturas murorum eius et ea quae corruerant instaurabo et reaedifi-

9:12 That they may possess the remnant of Edom, and of cabo eum sicut diebus antiquis 9:12 ut possideant reliquias

all the heathen, which are called by my name, saith the LORD that doeth this.

Idumeae et omnes nationes eo quod inuocatum sit nomen

9:13 Behold, the days come, saith the LORD, that the meum super eos dicit Dominus faciens haec 9:13 ecce dies

plowman shall overtake the reaper, and the treader of grapes him that soweth seed; and the mountains shall drop

sweet wine, and all the hills shall melt. ueniunt dicit Dominus et conprehendet arator messorem et

calcator uuae mittentem semen et stillabunt montes dulce-

9:14 And I will bring again the captivity of my people of dinem et omnes colles culti erunt 9:14 et conuertam capti-

Israel, and they shall build the waste cities, and inhabit them; and they shall plant vineyards, and drink the wine thereof; they shall also make gardens, and eat the fruit of them.

uitatem populi mei Israhel et aedificabunt ciuitates desertas et habitabunt et plantabunt uineas et bibent uinum earum et

9:15 And I will plant them upon their land, and they shall facient hortos et comedent fructus eorum 9:15 et plantabo

no more be pulled up out of their land which I have given them, saith the LORD thy God.

eos super humum suam et non euellam eos ultra de terra sua quam dedi eis dicit Dominus Deus tuus

Prophetia Abdiae

1:1 uisio Abdiae haec dicit Dominus Deus ad Edom audi- 1:1 The vision of Obadiah. Thus saith the Lord GOD con-

cerning Edom; We have heard a rumour from the LORD, and an ambassador is sent among the heathen, Arise ye,

tum audiuimus a Domino et legatum ad gentes misit surgite and let us rise up against her in battle. et consurgamus aduersum eum in proelium 1:2 ecce par- 1:2 Behold, I have made thee small among the heathen:

thou art greatly despised.

uulum te dedi in gentibus contemptibilis tu es ualde 1:3 su- 1:3 The pride of thine heart hath deceived thee, thou that

dwellest in the clefts of the rock, whose habitation is high; that saith in his heart, Who shall bring me down to the

perbia cordis tui extulit te habitantem in scissuris petrae ex- ground? altantem solium suum qui dicit in corde suo quis detrahet me in terram 1:4 si exaltatus fueris ut aquila et si inter sidera 1:4 Though thou exalt thyself as the eagle, and though

thou set thy nest among the stars, thence will I bring thee down, saith the LORD.

posueris nidum tuum inde detraham te dicit Dominus 1:5 si

1:5 If thieves came to thee, if robbers by night, (how art thou cut off!) would they not have stolen till they had enough? if the grapegatherers came to thee, would they not leave some grapes?

fures introissent ad te si latrones per noctem quomodo conticuisses nonne furati essent sufficientia sibi si uindemiatores introissent ad te numquid saltim racemos reliquissent tibi 1:6 quomodo scrutati sunt Esau inuestigauerunt abscondita 1:6 How are the things of Esau searched out! how are his

hidden things sought up!

eius 1:7 usque ad terminum emiserunt te omnes uiri foe- 1:7 All the men of thy confederacy have brought thee even to the border: the men that were at peace with thee have deceived thee, and prevailed against thee; that they eat thy bread have laid a wound under thee: there is none understanding in him.

deris tui inluserunt tibi inualuerunt aduersum te uiri pacis tuae qui comedunt tecum ponent insidias subter te non est prudentia in eo 1:8 numquid non in die illa dicit Dominus 1:8 Shall I not in that day, saith the LORD, even destroy

the wise men out of Edom, and understanding out of the mount of Esau?

perdam sapientes de Idumea et prudentiam de monte Esau 1:9 et timebunt fortes tui a meridie ut intereat uir de monte 1:9 And thy mighty men, O Teman, shall be dismayed, to

the end that every one of the mount of Esau may be cut off by slaughter.

Esau 1:10 propter interfectionem et propter iniquitatem in

1:10 For thy violence against thy brother Jacob shame shall cover thee, and thou shalt be cut off for ever.

fratrem tuum Iacob operiet te confusio et peribis in aeternum 1:11 in die cum stares aduersus quando capiebant alieni ex- 1:11 In the day that thou stoodest on the other side, in

the day that the strangers carried away captive his forces, and foreigners entered into his gates, and cast lots upon

ercitum eius et extranei ingrediebantur portas eius et super Jerusalem, even thou wast as one of them. Hierusalem mittebant sortem tu quoque eras quasi unus ex eis 1:12 et non despicies in die fratris tui in die peregri- 1:12 But thou shouldest not have looked on the day of

thy brother in the day that he became a stranger; neither shouldest thou have rejoiced over the children of Judah in the day of their destruction; neither shouldest thou have spoken proudly in the day of distress.

nationis eius et non laetaberis super filios Iuda in die perditionis eorum et non magnificabis os tuum in die angustiae 1:13 neque ingredieris portam populi mei in die ruinae eorum 1:13 Thou shouldest not have entered into the gate of my

people in the day of their calamity; yea, thou shouldest not have looked on their affliction in the day of their calamity, nor have laid hands on their substance in the day of their calamity;

neque despicies et tu in malis eius in die uastitatis illius et non emitteris aduersum exercitum eius in die uastitatis illius 1:14 neque stabis in exitibus ut interficias eos qui fugerint et 1:14 Neither shouldest thou have stood in the crossway, to

cut off those of his that did escape; neither shouldest thou have delivered up those of his that did remain in the day

non concludes reliquos eius in die tribulationis 1:15 quo- of distress.

1:15 For the day of the LORD is near upon all the heathen: as thou hast done, it shall be done unto thee: thy reward shall return upon thine own head.

niam iuxta est dies Domini super omnes gentes sicut fecisti fiet tibi retributionem tuam conuertet in caput tuum 1:16 1:16 For as ye have drunk upon my holy mountain, so shall

all the heathen drink continually, yea, they shall drink, and they shall swallow down, and they shall be as though they had not been.

Biblia Sacra Vulgata

quomodo enim bibisti super montem sanctum meum bibent omnes gentes iugiter et bibent et absorbent et erunt quasi

1:17 But upon mount Zion shall be deliverance, and there non sint 1:17 et in monte Sion erit saluatio et erit sanctus

shall be holiness; and the house of Jacob shall possess their possessions.

et possidebit domus Iacob eos qui se possederant 1:18 et

1:18 And the house of Jacob shall be a fire, and the house of Joseph a flame, and the house of Esau for stubble, and they shall kindle in them, and devour them; and there shall not be any remaining of the house of Esau; for the LORD hath spoken it.

erit domus Iacob ignis et domus Ioseph flamma et domus Esau stipula et succendentur in eis et deuorabunt eos et non

1:19 And they of the south shall possess the mount of erunt reliquiae domus Esau quia Dominus locutus est 1:19

Esau; and they of the plain the Philistines: and they shall possess the fields of Ephraim, and the fields of Samaria:

and Benjamin shall possess Gilead. et hereditabunt hii qui ad austrum montem Esau et qui in

campestribus Philisthim et possidebunt regionem Ephraim

1:20 And the captivity of this host of the children of Israel et regionem Samariae et Beniamin possidebit Galaad 1:20 shall possess that of the Canaanites, even unto Zarephath; and the captivity of Jerusalem, which is in Sepharad, shall

possess the cities of the south. et transmigratio exercitus huius filiorum Israhel omnia Cha-

naneorum usque ad Saraptham et transmigratio Hierusalem

1:21 And saviours shall come up on mount Zion to judge quae in Bosforo est possidebit ciuitates austri 1:21 et as-

the mount of Esau; and the kingdom shall be the LORD’s.

cendent saluatores in montem Sion iudicare montem Esau et erit Domino regnum

Prophetia Ionae

1:1 et factum est uerbum Domini ad Ionam filium Am- 1:1 Now the word of the LORD came unto Jonah the son

of Amittai, saying,

athi dicens 1:2 surge uade in Nineuen ciuitatem grandem 1:2 Arise, go to Nineveh, that great city, and cry against it;

for their wickedness is come up before me.

et praedica in ea quia ascendit malitia eius coram me 1:3 et 1:3 But Jonah rose up to flee unto Tarshish from the pres-

ence of the LORD, and went down to Joppa; and he found a ship going to Tarshish: so he paid the fare thereof, and went down into it, to go with them unto Tarshish from the presence of the LORD.

surrexit Iona ut fugeret in Tharsis a facie Domini et descendit Ioppen et inuenit nauem euntem in Tharsis et dedit naulum eius et descendit in eam ut iret cum eis in Tharsis a facie Domini 1:4 Dominus autem misit uentum magnum in mari 1:4 But the LORD sent out a great wind into the sea, and

there was a mighty tempest in the sea, so that the ship was like to be broken.

et facta est tempestas magna in mari et nauis periclitabatur conteri 1:5 et timuerunt nautae et clamauerunt uiri ad deum 1:5 Then the mariners were afraid, and cried every man

unto his god, and cast forth the wares that were in the ship into the sea, to lighten it of them. But Jonah was gone down into the sides of the ship; and he lay, and was fast asleep.

suum et miserunt uasa quae erant in naui in mare ut adleuiaretur ab eis et Iona descendit ad interiora nauis et dormiebat sopore graui 1:6 et accessit ad eum gubernator et dixit ei 1:6 So the shipmaster came to him, and said unto him,

What meanest thou, O sleeper? arise, call upon thy God, if so be that God will think upon us, that we perish not.

quid tu sopore deprimeris surge inuoca Deum tuum si forte recogitet Deus de nobis et non pereamus 1:7 et dixit uir ad 1:7 And they said every one to his fellow, Come, and let

us cast lots, that we may know for whose cause this evil is upon us. So they cast lots, and the lot fell upon Jonah.

collegam suum uenite et mittamus sortes et sciamus quare hoc malum sit nobis et miserunt sortes et cecidit sors super Ionam 1:8 et dixerunt ad eum indica nobis cuius causa 1:8 Then said they unto him, Tell us, we pray thee, for

whose cause this evil is upon us; What is thine occupation? and whence comest thou? what is thy country? and of

malum istud sit nobis quod est opus tuum quae terra tua et what people art thou? quo uel ex quo populo es tu 1:9 et dixit ad eos Hebraeus 1:9 And he said unto them, I am an Hebrew; and I fear the

LORD, the God of heaven, which hath made the sea and the dry land.

ego sum et Dominum Deum caeli ego timeo qui fecit mare et aridam 1:10 et timuerunt uiri timore magno et dixerunt 1:10 Then were the men exceedingly afraid, and said unto

him. Why hast thou done this? For the men knew that he fled from the presence of the LORD, because he had told

ad eum quid hoc fecisti cognouerunt enim uiri quod a fa- them. cie Domini fugeret quia indicauerat eis 1:11 et dixerunt ad 1:11 Then said they unto him, What shall we do unto thee,

that the sea may be calm unto us? for the sea wrought, and was tempestuous.

eum quid faciemus tibi et cessabit mare a nobis quia mare ibat et intumescebat 1:12 et dixit ad eos tollite me et mit- 1:12 And he said unto them, Take me up, and cast me forth

into the sea; so shall the sea be calm unto you: for I know that for my sake this great tempest is upon you.

tite in mare et cessabit mare a uobis scio enim ego quoniam propter me tempestas grandis haec super uos 1:13 et remi- 1:13 Nevertheless the men rowed hard to bring it to the

land; but they could not: for the sea wrought, and was tempestuous against them.

gabant uiri ut reuerterentur ad aridam et non ualebant quia mare ibat et intumescebat super eos 1:14 et clamauerunt ad 1:14 Wherefore they cried unto the LORD, and said, We

beseech thee, O LORD, we beseech thee, let us not perish for this man’s life, and lay not upon us innocent blood: for

Dominum et dixerunt quaesumus Domine ne pereamus in thou, O LORD, hast done as it pleased thee. anima uiri istius et ne des super nos sanguinem innocentem quia tu Domine sicut uoluisti fecisti 1:15 et tulerunt Ionam 1:15 So they look up Jonah, and cast him forth into the

sea: and the sea ceased from her raging.

et miserunt in mare et stetit mare a feruore suo 1:16 et ti- 1:16 Then the men feared the LORD exceedingly, and of-

fered a sacrifice unto the LORD, and made vows.

Biblia Sacra Vulgata

muerunt uiri timore magno Dominum et immolauerunt hostias Domino et uouerunt uota

2:1 Now the LORD had prepared a great fish to swallow 2:1 et praeparauit Dominus piscem grandem ut degluttiret

up Jonah. And Jonah was in the belly of the fish three days and three nights.

Ionam et erat Iona in uentre piscis tribus diebus et tribus noc-

2:2 Then Jonah prayed unto the LORD his God out of the tibus 2:2 et orauit Iona ad Dominum Deum suum de utero

fish’s belly,

2:3 And said, I cried by reason of mine affliction unto the piscis 2:3 et dixit clamaui de tribulatione mea ad Dominum

LORD, and he heard me; out of the belly of hell cried I, and thou heardest my voice.

et exaudiuit me de uentre inferni clamaui et exaudisti uocem

2:4 For thou hadst cast me into the deep, in the midst of the meam 2:4 et proiecisti me in profundum in corde maris et seas; and the floods compassed me about: all thy billows and thy waves passed over me.

flumen circumdedit me omnes gurgites tui et fluctus tui su-

2:5 Then I said, I am cast out of thy sight; yet I will look per me transierunt 2:5 et ego dixi abiectus sum a conspectu

again toward thy holy temple.

oculorum tuorum uerumtamen rursus uidebo templum sanc-

2:6 The waters compassed me about, even to the soul: the tum tuum 2:6 circumdederunt me aquae usque ad animam depth closed me round about, the weeds were wrapped about my head.

abyssus uallauit me pelagus operuit caput meum 2:7 ad ex-

2:7 I went down to the bottoms of the mountains; the earth with her bars was about me for ever: yet hast thou brought up my life from corruption, O LORD my God.

trema montium descendi terrae uectes concluserunt me in aeternum et subleuabis de corruptione uitam meam Domine

2:8 When my soul fainted within me I remembered the Deus meus 2:8 cum angustiaretur in me anima mea Do-LORD: and my prayer came in unto thee, into thine holy temple.

mini recordatus sum ut ueniat ad te oratio mea ad templum

2:9 They that observe lying vanities forsake their own sanctum tuum 2:9 qui custodiunt uanitates frustra miseri-

mercy. 2:10 But I will sacrifice unto thee with the voice of thanks- cordiam suam derelinquunt 2:10 ego autem in uoce laudis giving; I will pay that that I have vowed. Salvation is of the LORD.

immolabo tibi quaecumque uoui reddam pro salute Domino

2:11 And the LORD spake unto the fish, and it vomited 2:11 et dixit Dominus pisci et euomuit Ionam in aridam

out Jonah upon the dry land.

3:1 And the word of the LORD came unto Jonah the sec- 3:1 et factum est uerbum Domini ad Ionam secundo dicens

ond time, saying, 3:2 Arise, go unto Nineveh, that great city, and preach unto 3:2 surge uade ad Nineuen ciuitatem magnam et praedica in it the preaching that I bid thee. 3:3 So Jonah arose, and went unto Nineveh, according to ea praedicationem quam ego loquor ad te 3:3 et surrexit the word of the LORD. Now Nineveh was an exceeding great city of three days’ journey.

Iona et abiit in Nineuen iuxta uerbum Domini et Nineue erat

3:4 And Jonah began to enter into the city a day’s journey, ciuitas magna Dei itinere dierum trium 3:4 et coepit Iona and he cried, and said, Yet forty days, and Nineveh shall be overthrown.

introire in ciuitatem itinere diei unius et clamauit et dixit

3:5 So the people of Nineveh believed God, and pro- adhuc quadraginta dies et Nineue subuertetur 3:5 et credi-

claimed a fast, and put on sackcloth, from the greatest of them even to the least of them.

derunt uiri nineuitae in Deo et praedicauerunt ieiunium et

3:6 For word came unto the king of Nineveh, and he arose uestiti sunt saccis a maiore usque ad minorem 3:6 et per-

from his throne, and he laid his robe from him, and covered him with sackcloth, and sat in ashes.

uenit uerbum ad regem Nineue et surrexit de solio suo et abiecit uestimentum suum a se et indutus est sacco et sedit

3:7 And he caused it to be proclaimed and published in cinere 3:7 et clamauit et dixit in Nineue ex ore regis et

through Nineveh by the decree of the king and his nobles, saying, Let neither man nor beast, herd nor flock, taste any thing: let them not feed, nor drink water:

Prophetia Ionae

principum eius dicens homines et iumenta et boues et pecora non gustent quicquam nec pascantur et aquam non bibant 3:8 et operiantur saccis homines et iumenta et clament 3:8 But let man and beast be covered with sackcloth, and

cry mightily unto God: yea, let them turn every one from his evil way, and from the violence that is in their hands.

ad Dominum in fortitudine et conuertatur uir a uia sua mala et ab iniquitate quae est in manibus eorum 3:9 quis scit si 3:9 Who can tell if God will turn and repent, and turn away

from his fierce anger, that we perish not?

conuertatur et ignoscat Deus et reuertatur a furore irae suae et non peribimus 3:10 et uidit Deus opera eorum quia con- 3:10 And God saw their works, that they turned from their

evil way; and God repented of the evil, that he had said that he would do unto them; and he did it not.

uersi sunt a uia sua mala et misertus est Deus super malitiam quam locutus fuerat ut faceret eis et non fecit

4:1 et adflictus est Iona adflictione magna et iratus est 4:2 4:1 But it displeased Jonah exceedingly, and he was very

angry.

4:2 And he prayed unto the LORD, and said, I pray thee, O LORD, was not this my saying, when I was yet in my country? Therefore I fled before unto Tarshish: for I knew that thou art a gracious God, and merciful, slow to anger, and of great kindness, and repentest thee of the evil.

et orauit ad Dominum et dixit obsecro Domine numquid non hoc est uerbum meum cum adhuc essem in terra mea propter hoc praeoccupaui ut fugerem in Tharsis scio enim quia tu Deus clemens et misericors es patiens et multae miserationis et ignoscens super malitia 4:3 et nunc Domine tolle 4:3 Therefore now, O LORD, take, I beseech thee, my life

from me; for it is better for me to die than to live.

quaeso animam meam a me quia melior est mihi mors quam uita 4:4 et dixit Dominus putasne bene irasceris tu 4:5 et 4:4 Then said the LORD, Doest thou well to be angry?

4:5 So Jonah went out of the city, and sat on the east side of the city, and there made him a booth, and sat under it in the shadow, till he might see what would become of the city.

egressus est Iona de ciuitate et sedit contra orientem ciuitatis et fecit sibimet ibi umbraculum et sedebat subter eum in umbra donec uideret quid accideret ciuitati 4:6 et praepara- 4:6 And the LORD God prepared a gourd, and made it to

come up over Jonah, that it might be a shadow over his head, to deliver him from his grief. So Jonah was exceed-

uit Dominus Deus hederam et ascendit super caput Ionae ut ing glad of the gourd. esset umbra super caput eius et protegeret eum laborauerat enim et laetatus est Iona super hedera laetitia magna 4:7 et 4:7 But God prepared a worm when the morning rose the

next day, and it smote the gourd that it withered.

parauit Deus uermem ascensu diluculo in crastinum et percussit hederam et exaruit 4:8 et cum ortus fuisset sol prae- 4:8 And it came to pass, when the sun did arise, that God

prepared a vehement east wind; and the sun beat upon the head of Jonah, that he fainted, and wished in himself to

cepit Dominus uento calido et urenti et percussit sol super die, and said, It is better for me to die than to live. caput Ionae et aestuabat et petiuit animae suae ut moreretur et dixit melius est mihi mori quam uiuere 4:9 et dixit Do- 4:9 And God said to Jonah, Doest thou well to be angry

for the gourd? And he said, I do well to be angry, even unto death.

minus ad Ionam putasne bene irasceris tu super hederam et dixit bene irascor ego usque ad mortem 4:10 et dixit Domi- 4:10 Then said the LORD, Thou hast had pity on the

gourd, for the which thou hast not laboured, neither madest it grow; which came up in a night, and perished

nus tu doles super hederam in qua non laborasti neque fecisti in a night: ut cresceret quae sub una nocte nata est et una nocte periit 4:11 et ego non parcam Nineue ciuitati magnae in qua sunt 4:11 And should not I spare Nineveh, that great city,

wherein are more then sixscore thousand persons that cannot discern between their right hand and their left hand; and also much cattle?

Biblia Sacra Vulgata

plus quam centum uiginti milia hominum qui nesciunt quid sit inter dexteram et sinistram suam et iumenta multa

Prophetia Michaeae

1:1 uerbum Domini quod factum est ad Micham Morast- 1:1 The word of the LORD that came to Micah the

Morasthite in the days of Jotham, Ahaz, and Hezekiah, kings of Judah, which he saw concerning Samaria and

hiten in diebus Ioatham Ahaz Ezechiae regum Iuda quod ui- Jerusalem. dit super Samariam et Hierusalem 1:2 audite populi omnes 1:2 Hear, all ye people; hearken, O earth, and all that

therein is: and let the Lord GOD be witness against you, the LORD from his holy temple.

et adtendat terra et plenitudo eius et sit Dominus Deus uobis in testem Dominus de templo sancto suo 1:3 quia ecce 1:3 For, behold, the LORD cometh forth out of his place,

and will come down, and tread upon the high places of the earth.

Dominus egreditur de loco suo et descendet et calcabit super excelsa terrae 1:4 et consumentur montes subtus eum 1:4 And the mountains shall be molten under him, and the

valleys shall be cleft, as wax before the fire, and as the waters that are poured down a steep place.

et ualles scindentur sicut cera a facie ignis sicut aquae quae decurrunt in praeceps 1:5 in scelere Iacob omne istud et in 1:5 For the transgression of Jacob is all this, and for the

sins of the house of Israel. What is the transgression of Jacob? is it not Samaria? and what are the high places of

peccatis domus Israhel quod scelus Iacob nonne Samaria et Judah? are they not Jerusalem? quae excelsa Iudae nonne Hierusalem 1:6 et ponam Sama- 1:6 Therefore I will make Samaria as an heap of the field,

and as plantings of a vineyard: and I will pour down the stones thereof into the valley, and I will discover the foun-

riam quasi aceruum lapidum in agro cum plantatur uinea et dations thereof. detraham in uallem lapides eius et fundamenta eius reuelabo 1:7 et omnia sculptilia eius concidentur et omnes mercedes 1:7 And all the graven images thereof shall be beaten to

pieces, and all the hires thereof shall be burned with the fire, and all the idols thereof will I lay desolate: for she gathered it of the hire of an harlot, and they shall return to the hire of an harlot.

eius conburentur igni et omnia idola eius ponam in perditionem quia de mercedibus meretricis congregata sunt et usque ad mercedem meretricis reuertentur 1:8 super hoc plangam 1:8 Therefore I will wail and howl, I will go stripped and

naked: I will make a wailing like the dragons, and mourning as the owls.

et ululabo uadam spoliatus et nudus faciam planctum uelut draconum et luctum quasi strutionum 1:9 quia desperata est 1:9 For her wound is incurable; for it is come unto Judah;

he is come unto the gate of my people, even to Jerusalem.

plaga eius quia uenit usque ad Iudam tetigit portam populi mei usque ad Hierusalem 1:10 in Geth nolite adnuntiare la- 1:10 Declare ye it not at Gath, weep ye not at all: in the

house of Aphrah roll thyself in the dust.

crimis ne ploretis in domo Pulueris puluere uos conspergite 1:11 et transite uobis habitatio Pulchra confusa ignominia 1:11 Pass ye away, thou inhabitant of Saphir, having thy

shame naked: the inhabitant of Zaanan came not forth in the mourning of Bethezel; he shall receive of you his

non est egressa quae habitat in Exitu planctum domus Vici- standing. nae accipiet ex uobis quae stetit sibimet 1:12 quia infirmata 1:12 For the inhabitant of Maroth waited carefully for

good: but evil came down from the LORD unto the gate of Jerusalem.

est in bonum quae habitat in Amaritudinibus quia descendit

malum a Domino in portam Hierusalem 1:13 tumultus qua- 1:13 O thou inhabitant of Lachish, bind the chariot to the swift beast: she is the beginning of the sin to the daughter of Zion: for the transgressions of Israel were found in thee.

drigae stuporis habitanti Lachis principium peccati est filiae Sion quia in te inuenta sunt scelera Israhel 1:14 propterea 1:14 Therefore shalt thou give presents to Moreshethgath:

the houses of Achzib shall be a lie to the kings of Israel.

dabit emissarios super hereditatem Geth domus Mendacii in deceptionem regibus Israhel 1:15 adhuc heredem adducam 1:15 Yet will I bring an heir unto thee, O inhabitant of

Mareshah: he shall come unto Adullam the glory of Israel.

tibi quae habitas in Maresa usque Adollam ueniet gloria Israhel 1:16 decaluare et tondere super filios deliciarum tuarum 1:16 Make thee bald, and poll thee for thy delicate chil-

dren; enlarge thy baldness as the eagle; for they are gone into captivity from thee.

Biblia Sacra Vulgata

dilata caluitium tuum sicut aquila quoniam captiui ducti sunt ex te

2:1 Woe to them that devise iniquity, and work evil upon 2:1 uae qui cogitatis inutile et operamini malum in cubi-

their beds! when the morning is light, they practise it, because it is in the power of their hand.

libus uestris in luce matutina faciunt illud quoniam contra

2:2 And they covet fields, and take them by violence; and Deum est manus eorum 2:2 et concupierunt agros et uiohouses, and take them away: so they oppress a man and his house, even a man and his heritage.

lenter tulerunt et domos rapuerunt et calumniabantur uirum

2:3 Therefore thus saith the LORD; Behold, against this et domum eius uirum et hereditatem eius 2:3 idcirco haec family do I devise an evil, from which ye shall not remove your necks; neither shall ye go haughtily: for this time is

evil. dicit Dominus ecce ego cogito super familiam istam malum

unde non auferetis colla uestra et non ambulabitis superbi

2:4 In that day shall one take up a parable against you, and quoniam tempus pessimum est 2:4 in die illa sumetur sulament with a doleful lamentation, and say, We be utterly spoiled: he hath changed the portion of my people: how hath he removed it from me! turning away he hath divided our fields.

per uos parabola et cantabitur canticum cum suauitate dicentium depopulatione uastati sumus pars populi mei commutata est quomodo recedet a me cum reuertatur qui regiones

2:5 Therefore thou shalt have none that shall cast a cord nostras diuidat 2:5 propter hoc non erit tibi mittens funicu-

by lot in the congregation of the LORD.

2:6 Prophesy ye not, say they to them that prophesy: they lum sortis in coetu Domini 2:6 ne loquamini loquentes non

shall not prophesy to them, that they shall not take shame.

2:7 O thou that art named the house of Jacob, is the spirit stillabit super istos non conprehendet confusio 2:7 dicit do-

of the LORD straitened? are these his doings? do not my words do good to him that walketh uprightly?

mus Iacob numquid adbreuiatus est spiritus Domini aut tales sunt cogitationes eius nonne uerba mea bona sunt cum eo qui

2:8 Even of late my people is risen up as an enemy: ye recte graditur 2:8 et e contrario populus meus in aduersa-

pull off the robe with the garment from them that pass by securely as men averse from war.

rium consurrexit desuper tunica pallium sustulistis eos qui

2:9 The women of my people have ye cast out from their transiebant simpliciter conuertistis in bellum 2:9 mulieres

pleasant houses; from their children have ye taken away my glory for ever.

populi mei eiecistis de domo deliciarum suarum a paruulis

2:10 Arise ye, and depart; for this is not your rest: be- earum tulistis laudem meam in perpetuum 2:10 surgite et

cause it is polluted, it shall destroy you, even with a sore destruction.

ite quia non habetis hic requiem propter inmunditiam eius

2:11 If a man walking in the spirit and falsehood do lie, corrumpetur putredine pessima 2:11 utinam non essem uir

saying, I will prophesy unto thee of wine and of strong drink; he shall even be the prophet of this people.

habens spiritum et mendacium potius loquerer stillabo tibi in uinum et in ebrietatem et erit super quem stillatur popu-

2:12 I will surely assemble, O Jacob, all of thee; I will lus iste 2:12 congregatione congregabo Iacob totum te in

surely gather the remnant of Israel; I will put them together as the sheep of Bozrah, as the flock in the midst of their fold: they shall make great noise by reason of the multitude of men.

unum conducam reliquias Israhel pariter ponam illum quasi gregem in ouili quasi pecus in medio caularum tumultuabun-

2:13 The breaker is come up before them: they have bro- tur a multitudine hominum 2:13 ascendet enim pandens iter

ken up, and have passed through the gate, and are gone out by it: and their king shall pass before them, and the LORD

on the head of them. ante eos diuident et transibunt portam et egredientur per eam et transibit rex eorum coram eis et Dominus in capite eorum

Prophetia Michaeae

3:1 et dixi audite principes Iacob et duces domus Israhel 3:1 And I said, Hear, I pray you, O heads of Jacob, and

ye princes of the house of Israel; Is it not for you to know judgment?

numquid non uestrum est scire iudicium 3:2 qui odio ha-

3:2 Who hate the good, and love the evil; who pluck off their skin from off them, and their flesh from off their bones;

betis bonum et diligitis malum qui uiolenter tollitis pelles eorum desuper eos et carnem eorum desuper ossibus eorum 3:3 qui comederunt carnem populi mei et pellem eorum de- 3:3 Who also eat the flesh of my people, and flay their skin

from off them; and they break their bones, and chop them in pieces, as for the pot, and as flesh within the caldron.

super excoriauerunt et ossa eorum confregerunt et conciderunt sicut in lebete et quasi carnem in medio ollae 3:4 tunc 3:4 Then shall they cry unto the LORD, but he will not

hear them: he will even hide his face from them at that time, as they have behaved themselves ill in their doings.

clamabunt ad Dominum et non exaudiet eos et abscondet faciem suam ab eis in tempore illo sicut nequiter egerunt in adinuentionibus suis 3:5 haec dicit Dominus super pro- 3:5 Thus saith the LORD concerning the prophets that

make my people err, that bite with their teeth, and cry, Peace; and he that putteth not into their mouths, they even

phetas qui seducunt populum meum qui mordent dentibus prepare war against him. suis et praedicant pacem et si quis non dederit in ore eorum quippiam sanctificant super eum proelium 3:6 prop- 3:6 Therefore night shall be unto you, that ye shall not

have a vision; and it shall be dark unto you, that ye shall not divine; and the sun shall go down over the prophets,

terea nox uobis pro uisione erit et tenebrae uobis pro diui- and the day shall be dark over them. natione et occumbet sol super prophetas et obtenebrabitur super eos dies 3:7 et confundentur qui uident uisiones et 3:7 Then shall the seers be ashamed, and the diviners con-

founded: yea, they shall all cover their lips; for there is no answer of God.

confundentur diuini et operient uultus suos omnes quia non est responsum Dei 3:8 uerumtamen ego repletus sum for- 3:8 But truly I am full of power by the spirit of the LORD,

and of judgment, and of might, to declare unto Jacob his transgression, and to Israel his sin.

titudine spiritus Domini iudicio et uirtute ut adnuntiem Iacob scelus suum et Israhel peccatum suum 3:9 audite haec 3:9 Hear this, I pray you, ye heads of the house of Jacob,

and princes of the house of Israel, that abhor judgment, and pervert all equity.

principes domus Iacob et iudices domus Israhel qui abominamini iudicium et omnia recta peruertitis 3:10 qui aedifi- 3:10 They build up Zion with blood, and Jerusalem with

iniquity.

catis Sion in sanguinibus et Hierusalem in iniquitate 3:11 3:11 The heads thereof judge for reward, and the priests

thereof teach for hire, and the prophets thereof divine for money: yet will they lean upon the LORD, and say, Is not

principes eius in muneribus iudicabant et sacerdotes eius in the LORD among us? none evil can come upon us. mercede docebant et prophetae eius in pecunia diuinabant et super Dominum requiescebant dicentes numquid non Dominus in medio nostrum non uenient super nos mala 3:12 3:12 Therefore shall Zion for your sake be plowed as a

field, and Jerusalem shall become heaps, and the mountain of the house as the high places of the forest.

propter hoc causa uestri Sion quasi ager arabitur et Hierusalem quasi aceruus lapidum erit et mons templi in excelsa siluarum

4:1 et in nouissimo dierum erit mons domus Domini prae- 4:1 But in the last days it shall come to pass, that the

mountain of the house of the LORD shall be established in the top of the mountains, and it shall be exalted above

paratus in uertice montium et sublimis super colles et fluent the hills; and people shall flow unto it. ad eum populi 4:2 et properabunt gentes multae et dicent 4:2 And many nations shall come, and say, Come, and let

us go up to the mountain of the LORD, and to the house of the God of Jacob; and he will teach us of his ways, and we will walk in his paths: for the law shall go forth of Zion, and the word of the LORD from Jerusalem.

Biblia Sacra Vulgata

uenite ascendamus ad montem Domini et ad domum Dei Iacob et docebit nos de uiis suis et ibimus in semitis eius quia de Sion egredietur lex et uerbum Domini de Hierusa-

4:3 And he shall judge among many people, and rebuke lem 4:3 et iudicabit inter populos multos et corripiet gen-

strong nations afar off; and they shall beat their swords into plowshares, and their spears into pruninghooks: nation shall not lift up a sword against nation, neither shall they learn war any more.

tes fortes usque in longinquum et concident gladios suos in uomeres et hastas suas in ligones non sumet gens aduersus

4:4 But they shall sit every man under his vine and un- gentem gladium et non discent ultra belligerare 4:4 et sede-

der his fig tree; and none shall make them afraid: for the mouth of the LORD of hosts hath spoken it.

bit uir subtus uineam suam et subtus ficum suam et non erit

4:5 For all people will walk every one in the name of his qui deterreat quia os Domini exercituum locutum est 4:5

god, and we will walk in the name of the LORD our God for ever and ever.

quia omnes populi ambulabunt unusquisque in nomine dei sui nos autem ambulabimus in nomine Domini Dei nostri

4:6 In that day, saith the LORD, will I assemble her that in aeternum et ultra 4:6 in die illa dicit Dominus congre-

halteth, and I will gather her that is driven out, and her that I have afflicted;

gabo claudicantem et eam quam eieceram colligam et quam

4:7 And I will make her that halted a remnant, and her that adflixeram 4:7 et ponam claudicantem in reliquias et eam was cast far off a strong nation: and the LORD shall reign over them in mount Zion from henceforth, even for ever.

quae laborauerat in gentem robustam et regnabit Dominus super eos in monte Sion ex hoc nunc et usque in aeternum

4:8 And thou, O tower of the flock, the strong hold of 4:8 et tu turris Gregis nebulosa filiae Sion usque ad te ueniet

the daughter of Zion, unto thee shall it come, even the first dominion; the kingdom shall come to the daughter of

Jerusalem. et ueniet potestas prima regnum filiae Hierusalem 4:9 nunc

4:9 Now why dost thou cry out aloud? is there no king in thee? is thy counsellor perished? for pangs have taken thee as a woman in travail.

quare maerore contraheris numquid rex non est tibi aut consiliarius tuus periit quia conprehendit te dolor sicut partu-

4:10 Be in pain, and labour to bring forth, O daughter of rientem 4:10 dole et satage filia Sion quasi parturiens quia

Zion, like a woman in travail: for now shalt thou go forth out of the city, and thou shalt dwell in the field, and thou shalt go even to Babylon; there shalt thou be delivered; there the LORD shall redeem thee from the hand of thine enemies.

nunc egredieris de ciuitate et habitabis in regione et uenies usque ad Babylonem ibi liberaberis ibi rediiet te Dominus

4:11 Now also many nations are gathered against thee, that de manu inimicorum tuorum 4:11 et nunc congregatae sunt say, Let her be defiled, and let our eye look upon Zion.

super te gentes multae quae dicunt lapidetur et aspiciat in

4:12 But they know not the thoughts of the LORD, neither Sion oculus noster 4:12 ipsi autem non cognouerunt cogita-

understand they his counsel: for he shall gather them as the sheaves into the floor.

tiones Domini et non intellexerunt consilium eius quia con-

4:13 Arise and thresh, O daughter of Zion: for I will make gregauit eos quasi faenum areae 4:13 surge et tritura filia

thine horn iron, and I will make thy hoofs brass: and thou shalt beat in pieces many people: and I will consecrate their gain unto the LORD, and their substance unto the Lord of the whole earth.

Sion quia cornu tuum ponam ferreum et ungulas tuas ponam aereas et comminues populos multos et interficiam Domino rapinas eorum et fortitudinem eorum Domino uniuersae terrae

5:1 Now gather thyself in troops, O daughter of troops: he 5:1 nunc uastaberis filia latronis obsidionem posuerunt su-

hath laid siege against us: they shall smite the judge of Israel with a rod upon the cheek.

Prophetia Michaeae

per nos in uirga percutient maxillam iudicis Israhel 5:2 et 5:2 But thou, Bethlehem Ephratah, though thou be little

among the thousands of Judah, yet out of thee shall he come forth unto me that is to be ruler in Israel; whose

tu Bethleem Ephrata paruulus es in milibus Iuda ex te mihi goings forth have been from of old, from everlasting. egredietur qui sit dominator in Israhel et egressus eius ab initio a diebus aeternitatis 5:3 propter hoc dabit eos usque ad 5:3 Therefore will he give them up, until the time that she

which travaileth hath brought forth: then the remnant of his brethren shall return unto the children of Israel.

tempus in quo parturiens pariet reliquiae fratrum eius conuertentur ad filios Israhel 5:4 et stabit et pascet in fortitu- 5:4 And he shall stand and feed in the strength of the

LORD, in the majesty of the name of the LORD his God; and they shall abide: for now shall he be great unto the

dine Domini in sublimitate nominis Domini Dei sui et con- ends of the earth. uertentur quia nunc magnificabitur usque ad terminos terrae 5:5 et erit iste pax Assyrius cum uenerit in terram nostram 5:5 And this man shall be the peace, when the Assyrian

shall come into our land: and when he shall tread in our palaces, then shall we raise against him seven shepherds,

et quando calcauerit in domibus nostris et suscitabimus su- and eight principal men. per eum septem pastores et octo primates homines 5:6 et 5:6 And they shall waste the land of Assyria with the

sword, and the land of Nimrod in the entrances thereof: thus shall he deliver us from the Assyrian, when he cometh

pascent terram Assur in gladio et terram Nemrod in lanceis into our land, and when he treadeth within our borders. eius et liberabit ab Assur cum uenerit in terram nostram et cum calcauerit in finibus nostris 5:7 et erunt reliquiae Ia- 5:7 And the remnant of Jacob shall be in the midst of many

people as a dew from the LORD, as the showers upon the grass, that tarrieth not for man, nor waiteth for the sons of

cob in medio populorum multorum quasi ros a Domino et men. quasi stillae super herbam quae non expectat uirum et non praestolatur filios hominum 5:8 et erunt reliquiae Iacob in 5:8 And the remnant of Jacob shall be among the Gentiles

in the midst of many people as a lion among the beasts of the forest, as a young lion among the flocks of sheep: who, if he go through, both treadeth down, and teareth in pieces, and none can deliver.

gentibus in medio populorum multorum quasi leo in iumentis siluarum et quasi catulus leonis in gregibus pecorum qui cum transierit et conculcauerit et ceperit non est qui eruat 5:9 exaltabitur manus tua super hostes tuos et omnes inimici 5:9 Thine hand shall be lifted up upon thine adversaries,

and all thine enemies shall be cut off.

tui interibunt 5:10 et erit in die illa dicit Dominus auferam 5:10 And it shall come to pass in that day, saith the LORD,

that I will cut off thy horses out of the midst of thee, and I will destroy thy chariots:

equos tuos de medio tui et disperdam quadrigas tuas 5:11 et

5:11 And I will cut off the cities of thy land, and throw down all thy strong holds: And I will cut off witchcrafts out of thine hand; and thou shalt have no more soothsayers:

perdam ciuitates terrae tuae et destruam omnes munitiones tuas et auferam maleficia de manu tua et diuinationes non erunt in te 5:12 et perire faciam sculptilia tua et statuas tuas 5:12 Thy graven images also will I cut off, and thy stand-

ing images out of the midst of thee; and thou shalt no more worship the work of thine hands.

de medio tui et non adorabis ultra opera manuum tuarum 5:13 et euellam lucos tuos de medio tui et conteram ciuitates 5:13 And I will pluck up thy groves out of the midst of

thee: so will I destroy thy cities.

tuas 5:14 et faciam in furore et in indignatione ultionem in 5:14 And I will execute vengeance in anger and fury upon

the heathen, such as they have not heard.

omnibus gentibus quae non audierunt

6:1 audite quae Dominus loquitur surge contende iudicio 6:1 Hear ye now what the LORD saith; Arise, contend

thou before the mountains, and let the hills hear thy voice.

aduersum montes et audiant colles uocem tuam 6:2 audiant 6:2 Hear ye, O mountains, the LORD’s controversy, and

ye strong foundations of the earth: for the LORD hath a controversy with his people, and he will plead with Israel.

montes iudicium Domini et fortia fundamenta terrae quia

Biblia Sacra Vulgata

iudicium Domini cum populo suo et cum Israhel diiudica-

6:3 O my people, what have I done unto thee? and wherein bitur 6:3 populus meus quid feci tibi et quid molestus fui

have I wearied thee? testify against me.

6:4 For I brought thee up out of the land of Egypt, and re- tibi responde mihi 6:4 quia eduxi te de terra Aegypti et de

deemed thee out of the house of servants; and I sent before thee Moses, Aaron, and Miriam.

domo seruientium liberaui te et misi ante faciem tuam Mo-

6:5 O my people, remember now what Balak king of Moab sen et Aaron et Mariam 6:5 populus meus memento quaeso

consulted, and what Balaam the son of Beor answered him from Shittim unto Gilgal; that ye may know the righteous-

ness of the LORD. quid cogitauerit Balac rex Moab et quid responderit ei Ba-

laam filius Beor de Setthim usque ad Galgalam ut cognosce-

6:6 Wherewith shall I come before the LORD, and bow ret iustitias Domini 6:6 quid dignum offeram Domino cur-

myself before the high God? shall I come before him with burnt offerings, with calves of a year old?

uem genu Deo excelso numquid offeram ei holocaustomata

6:7 Will the LORD be pleased with thousands of rams, et uitulos anniculos 6:7 numquid placari potest Dominus

or with ten thousands of rivers of oil? shall I give my firstborn for my transgression, the fruit of my body for the

sin of my soul? in milibus arietum aut in multis milibus hircorum pinguium

numquid dabo primogenitum meum pro scelere meo fruc-

6:8 He hath shewed thee, O man, what is good; and what tum uentris mei pro peccato animae meae 6:8 indicabo tibi

doth the LORD require of thee, but to do justly, and to love mercy, and to walk humbly with thy God?

o homo quid sit bonum et quid Dominus quaerat a te utique facere iudicium et diligere misericordiam et sollicitum am-

6:9 The LORD’s voice crieth unto the city, and the man of bulare cum Deo tuo 6:9 uox Domini ad ciuitatem clamat et wisdom shall see thy name: hear ye the rod, and who hath appointed it.

salus erit timentibus nomen tuum audite tribus et quis adpro-

6:10 Are there yet the treasures of wickedness in the house babit illud 6:10 adhuc ignis in domo impii thesauri iniqui-

of the wicked, and the scant measure that is abominable?

6:11 Shall I count them pure with the wicked balances, tatis et mensura minor irae plena 6:11 numquid iustificabo

and with the bag of deceitful weights?

6:12 For the rich men thereof are full of violence, and the stateram impiam et saccelli pondera dolosa 6:12 in quibus

inhabitants thereof have spoken lies, and their tongue is deceitful in their mouth.

diuites eius repleti sunt iniquitate et habitantes in ea loquebantur mendacium et lingua eorum fraudulenta in ore eorum

6:13 Therefore also will I make thee sick in smiting thee, 6:13 et ego ergo coepi percutere te perditione super peccatis

in making thee desolate because of thy sins.

6:14 Thou shalt eat, but not be satisfied; and thy casting tuis 6:14 tu comedes et non saturaberis et humiliatio tua in

down shall be in the midst of thee; and thou shalt take hold, but shalt not deliver; and that which thou deliverest

will I give up to the sword. medio tui et adprehendes et non saluabis et quos saluaueris 6:15 Thou shalt sow, but thou shalt not reap; thou shalt in gladium dabo 6:15 tu seminabis et non metes tu calcabis

tread the olives, but thou shalt not anoint thee with oil; and sweet wine, but shalt not drink wine.

oliuam et non ungueris oleo et mustum et non bibes uinum

6:16 For the statutes of Omri are kept, and all the works 6:16 et custodisti praecepta Omri et omne opus domus Achab

of the house of Ahab, and ye walk in their counsels; that I should make thee a desolation, and the inhabitants thereof an hissing: therefore ye shall bear the reproach of my people.

et ambulasti in uoluntatibus eorum ut darem te in perditionem et habitantes in ea in sibilum et obprobrium populi mei portabitis

7:1 Woe is me! for I am as when they have gathered the 7:1 uae mihi quia factus sum sicut qui colligit in autumno

summer fruits, as the grapegleanings of the vintage: there is no cluster to eat: my soul desired the firstripe fruit.

racemos uindemiae non est botrus ad comedendum praeco-

Prophetia Michaeae

quas ficus desiderauit anima mea 7:2 periit sanctus de terra 7:2 The good man is perished out of the earth: and there

is none upright among men: they all lie in wait for blood; they hunt every man his brother with a net.

et rectus in hominibus non est omnes in sanguine insidiantur

uir fratrem suum uenatur ad mortem 7:3 malum manuum 7:3 That they may do evil with both hands earnestly, the prince asketh, and the judge asketh for a reward; and the great man, he uttereth his mischievous desire: so they

suarum dicunt bonum princeps postulat et iudex in reddendo wrap it up. est et magnus locutus est desiderium animae suae et conturbauerunt eam 7:4 qui optimus in eis est quasi paliurus et 7:4 The best of them is as a brier: the most upright is

sharper than a thorn hedge: the day of thy watchmen and thy visitation cometh; now shall be their perplexity.

qui rectus quasi spina de sepe dies speculationis tuae uisitatio tua uenit nunc erit uastitas eorum 7:5 nolite credere 7:5 Trust ye not in a friend, put ye not confidence in a

guide: keep the doors of thy mouth from her that lieth in thy bosom.

amico et nolite confidere in duce ab ea quae dormit in sinu tuo custodi claustra oris tui 7:6 quia filius contumeliam fa- 7:6 For the son dishonoureth the father, the daughter riseth

up against her mother, the daughter in law against her mother in law; a man’s enemies are the men of his own

cit patri filia consurgit aduersus matrem suam nurus contra house. socrum suam inimici hominis domestici eius 7:7 ego autem 7:7 Therefore I will look unto the LORD; I will wait for

the God of my salvation: my God will hear me.

ad Dominum aspiciam expectabo Deum saluatorem meum audiet me Deus meus 7:8 ne laeteris inimica mea super me 7:8 Rejoice not against me, O mine enemy: when I fall,

I shall arise; when I sit in darkness, the LORD shall be a light unto me.

quia cecidi consurgam cum sedero in tenebris Dominus lux mea est 7:9 iram Domini portabo quoniam peccaui ei do- 7:9 I will bear the indignation of the LORD, because I

have sinned against him, until he plead my cause, and execute judgment for me: he will bring me forth to the light,

nec iudicet causam meam et faciat iudicium meum educet and I shall behold his righteousness. me in lucem uidebo in iustitiam eius 7:10 et aspiciet in- 7:10 Then she that is mine enemy shall see it, and shame

shall cover her which said unto me, Where is the LORD thy God? mine eyes shall behold her: now shall she be

imica mea et operietur confusione quae dicit ad me ubi est trodden down as the mire of the streets. Dominus Deus tuus oculi mei uidebunt in eam nunc erit in conculcationem ut lutum platearum 7:11 dies ut aedificen- 7:11 In the day that thy walls are to be built, in that day

shall the decree be far removed.

tur maceriae tuae in die illa longe fiet lex 7:12 in die illa et 7:12 In that day also he shall come even to thee from As-

syria, and from the fortified cities, and from the fortress even to the river, and from sea to sea, and from mountain

usque ad te ueniet Assur et usque ad ciuitates munitas et a to mountain. ciuitatibus munitis usque ad flumen et ad mare de mari et ad

montem de monte 7:13 et erit terra in desolationem prop- 7:13 Notwithstanding the land shall be desolate because

of them that dwell therein, for the fruit of their doings.

ter habitatores suos et propter fructum cogitationum eorum 7:14 pasce populum tuum in uirga tua gregem hereditatis tuae 7:14 Feed thy people with thy rod, the flock of thine her-

itage, which dwell solitarily in the wood, in the midst of Carmel: let them feed in Bashan and Gilead, as in the days

habitantes solos in saltu in medio Carmeli pascentur Basan of old. et Galaad iuxta dies antiquos 7:15 secundum dies egressio- 7:15 According to the days of thy coming out of the land

of Egypt will I shew unto him marvellous things.

nis tuae de terra Aegypti ostendam ei mirabilia 7:16 uide- 7:16 The nations shall see and be confounded at all their

might: they shall lay their hand upon their mouth, their ears shall be deaf.

bunt gentes et confundentur super omni fortitudine sua po-

nent manus super os aures eorum surdae erunt 7:17 lingent 7:17 They shall lick the dust like a serpent, they shall move out of their holes like worms of the earth: they shall be afraid of the LORD our God, and shall fear because of

puluerem sicut serpens uelut reptilia terrae proturbabuntur thee.

Biblia Sacra Vulgata

de aedibus suis Dominum Deum nostrum desiderabunt et ti-

7:18 Who is a God like unto thee, that pardoneth iniq- mebunt te 7:18 quis Deus similis tui qui aufers iniquitatem

uity, and passeth by the transgression of the remnant of his heritage? he retaineth not his anger for ever, because

he delighteth in mercy. et transis peccatum reliquiarum hereditatis tuae non inmittet 7:19 He will turn again, he will have compassion upon us; ultra furorem suum quoniam uolens misericordiam est 7:19

he will subdue our iniquities; and thou wilt cast all their sins into the depths of the sea.

reuertetur et miserebitur nostri deponet iniquitates nostras

7:20 Thou wilt perform the truth to Jacob, and the mercy et proiciet in profundum maris omnia peccata nostra 7:20

to Abraham, which thou hast sworn unto our fathers from the days of old.

dabis ueritatem Iacob misericordiam Abraham quae iurasti patribus nostris a diebus antiquis

Prophetia Nahum

1:1 onus Nineue liber uisionis Naum Helcesei 1:2 Deus 1:1 The burden of Nineveh. The book of the vision of

Nahum the Elkoshite.

1:2 God is jealous, and the LORD revengeth; the LORD revengeth, and is furious; the LORD will take vengeance on his adversaries, and he reserveth wrath for his enemies.

aemulator et ulciscens Dominus ulciscens Dominus et habens furorem ulciscens Dominus in hostes suos et irascens ipse inimicis suis 1:3 Dominus patiens et magnus fortitu- 1:3 The LORD is slow to anger, and great in power, and

will not at all acquit the wicked: the LORD hath his way in the whirlwind and in the storm, and the clouds are the

dine et mundans non faciet innocentem Dominus in tempe- dust of his feet. state et turbine uiae eius et nebulae puluis pedum eius 1:4 1:4 He rebuketh the sea, and maketh it dry, and drieth up

all the rivers: Bashan languisheth, and Carmel, and the flower of Lebanon languisheth.

increpans mare et exsiccans illud et omnia flumina ad desertum deducens infirmatus est Basan et Carmelus et flos Libani elanguit 1:5 montes commoti sunt ab eo et colles 1:5 The mountains quake at him, and the hills melt, and

the earth is burned at his presence, yea, the world, and all that dwell therein.

adsolati sunt et contremuit terra a facie eius et orbis et omnes habitantes in eo 1:6 ante faciem indignationis eius quis 1:6 Who can stand before his indignation? and who can

abide in the fierceness of his anger? his fury is poured out like fire, and the rocks are thrown down by him.

stabit et quis resistet in ira furoris eius indignatio eius effusa est ut ignis et petrae dissolutae sunt ab eo 1:7 bonus Do- 1:7 The LORD is good, a strong hold in the day of trouble;

and he knoweth them that trust in him.

minus et confortans in die tribulationis et sciens sperantes in se 1:8 et in diluuio praetereunte consummationem faciet 1:8 But with an overrunning flood he will make an utter

end of the place thereof, and darkness shall pursue his enemies.

loci eius et inimicos eius persequentur tenebrae 1:9 quid

1:9 What do ye imagine against the LORD? he will make an utter end: affliction shall not rise up the second time.

cogitatis contra Dominum consummationem ipse faciet non consurget duplex tribulatio 1:10 quia sicut spinae se inui- 1:10 For while they be folden together as thorns, and while

they are drunken as drunkards, they shall be devoured as stubble fully dry.

cem conplectuntur sic conuiuium eorum pariter potantium consumentur quasi stipula ariditate plena 1:11 ex te exiuit 1:11 There is one come out of thee, that imagineth evil

against the LORD, a wicked counsellor.

cogitans contra Dominum malitiam mente pertractans praeuaricationem 1:12 haec dicit Dominus si perfecti fuerint et 1:12 Thus saith the LORD; Though they be quiet, and like-

wise many, yet thus shall they be cut down, when he shall pass through. Though I have afflicted thee, I will afflict

ita plures sic quoque adtondentur et pertransibit adflixi te et thee no more. non adfligam te ultra 1:13 et nunc conteram uirgam eius de 1:13 For now will I break his yoke from off thee, and will

burst thy bonds in sunder.

dorso tuo et uincula tua disrumpam 1:14 et praecipiet su- 1:14 And the LORD hath given a commandment concern-

ing thee, that no more of thy name be sown: out of the house of thy gods will I cut off the graven image and the

per te Dominus non seminabitur ex nomine tuo amplius de molten image: I will make thy grave; for thou art vile. domo Dei tui interficiam sculptile et conflatile ponam sepulchrum tuum quia inhonoratus es 1:15 ecce super mon- 1:15 Behold upon the mountains the feet of him that

bringeth good tidings, that publisheth peace! O Judah, keep thy solemn feasts, perform thy vows: for the wicked

tes pedes euangelizantis et adnuntiantis pacem celebra Iuda shall no more pass through thee; he is utterly cut off. festiuitates tuas et redde uota tua quia non adiciet ultra ut pertranseat in te Belial uniuersus interiit

2:1 ascendit qui dispergat coram te qui custodit obsi- 2:1 He that dasheth in pieces is come up before thy face:

keep the munition, watch the way, make thy loins strong, fortify thy power mightily.

dionem contemplare uiam conforta lumbos robora uirtutem

Biblia Sacra Vulgata

2:2 For the LORD hath turned away the excellency of Ja- ualde 2:2 quia reddidit Dominus superbiam Iacob sicut su-

cob, as the excellency of Israel: for the emptiers have emptied them out, and marred their vine branches.

perbiam Israhel quia uastatores dissipauerunt eos et propa-

2:3 The shield of his mighty men is made red, the valiant gines eorum corruperunt 2:3 clypeus fortium eius ignitus

men are in scarlet: the chariots shall be with flaming torches in the day of his preparation, and the fir trees shall

be terribly shaken. uiri exercitus in coccineis igneae habenae currus in die prae-2:4 The chariots shall rage in the streets, they shall justle parationis eius et agitatores consopiti sunt 2:4 in itineribus

one against another in the broad ways: they shall seem like torches, they shall run like the lightnings.

conturbati sunt quadrigae conlisae sunt in plateis aspectus

2:5 He shall recount his worthies: they shall stumble in eorum quasi lampades quasi fulgura discurrentia 2:5 recor-

their walk; they shall make haste to the wall thereof, and the defence shall be prepared.

dabitur fortium suorum ruent in itineribus suis uelociter as-

2:6 The gates of the rivers shall be opened, and the palace cendent muros eius et praeparabitur umbraculum 2:6 porshall be dissolved.

tae fluuiorum apertae sunt et templum ad solum dirutum

2:7 And Huzzab shall be led away captive, she shall be 2:7 et miles captiuus abductus est et ancillae eius minabantur

brought up, and her maids shall lead her as with the voice of doves, tabering upon their breasts.

gementes ut columbae murmurantes in cordibus suis 2:8 et

2:8 But Nineveh is of old like a pool of water: yet they shall flee away. Stand, stand, shall they cry; but none shall look back.

Nineue quasi piscina aquarum aquae eius ipsi uero fugerunt

2:9 Take ye the spoil of silver, take the spoil of gold: for state state et non est qui reuertatur 2:9 diripite argentum

there is none end of the store and glory out of all the pleasant furniture.

diripite aurum et non est finis diuitiarum ex omnibus uasis

2:10 She is empty, and void, and waste: and the heart desiderabilibus 2:10 dissipata et scissa et dilacerata et cor

melteth, and the knees smite together, and much pain is in all loins, and the faces of them all gather blackness.

tabescens et dissolutio geniculorum et defectio in cunctis re-

2:11 Where is the dwelling of the lions, and the feeding- nibus et facies omnium sicut nigredo ollae 2:11 ubi est ha-

place of the young lions, where the lion, even the old lion, walked, and the lion’s whelp, and none made them afraid?

bitaculum leonum et pascua catulorum leonum ad quam iuit leo ut ingrederetur illuc catulus leonis et non est qui exter-

2:12 The lion did tear in pieces enough for his whelps, and reat 2:12 leo cepit sufficienter catulis suis et necauit leaenis

strangled for his lionesses, and filled his holes with prey, and his dens with ravin.

suis et impleuit praeda speluncas suas et cubile suum rapina

2:13 Behold, I am against thee, saith the LORD of hosts, 2:13 ecce ego ad te dicit Dominus exercituum et succendam

and I will burn her chariots in the smoke, and the sword shall devour thy young lions: and I will cut off thy prey from the earth, and the voice of thy messengers shall no more be heard.

usque ad fumum quadrigas eius et leunculos tuos comedet gladius et exterminabo de terra praedam tuam et non audietur ultra uox nuntiorum tuorum

3:1 Woe to the bloody city! it is all full of lies and robbery; 3:1 uae ciuitas sanguinum uniuersa mendacii dilaceratione

the prey departeth not;

3:2 The noise of a whip, and the noise of the rattling of plena non recedet a te rapina 3:2 uox flagelli et uox impetus

the wheels, and of the pransing horses, and of the jumping chariots.

rotae et equi frementis et quadrigae feruentis equitis ascen-

3:3 The horseman lifteth up both the bright sword and the dentis 3:3 et micantis gladii et fulgurantis hastae et multi-

glittering spear: and there is a multitude of slain, and a great number of carcases; and there is none end of their

corpses; they stumble upon their corpses: tudinis interfectae et grauis ruinae nec est finis cadauerum et 3:4 Because of the multitude of the whoredoms of the corruent in corporibus suis 3:4 propter multitudinem forni-

wellfavoured harlot, the mistress of witchcrafts, that selleth nations through her whoredoms, and families through

her witchcrafts. cationum meretricis speciosae et gratae et habentis malefi-

Prophetia Nahum

cia quae uendidit gentes in fornicationibus suis et familias in maleficiis suis 3:5 ecce ego ad te dicit Dominus exercituum 3:5 Behold, I am against thee, saith the LORD of hosts;

and I will discover thy skirts upon thy face, and I will shew the nations thy nakedness, and the kingdoms thy shame.

et reuelabo pudenda tua in facie tua et ostendam gentibus nuditatem tuam et regnis ignominiam tuam 3:6 et proiciam 3:6 And I will cast abominable filth upon thee, and make

thee vile, and will set thee as a gazingstock.

super te abominationes et contumeliis te adficiam et ponam te in exemplum 3:7 et erit omnis qui uiderit te resiliet a 3:7 And it shall come to pass, that all they that look upon

thee shall flee from thee, and say, Nineveh is laid waste: who will bemoan her? whence shall I seek comforters for

te et dicet uastata est Nineue quis commouebit super te ca- thee? put unde quaeram consolatorem tibi 3:8 numquid melior es 3:8 Art thou better than populous No, that was situate

among the rivers, that had the waters round about it, whose rampart was the sea, and her wall was from the sea?

ab Alexandria populorum quae habitat in fluminibus aqua in circuitu eius cuius diuitiae mare aquae muri eius 3:9 Aet- 3:9 Ethiopia and Egypt were her strength, and it was infi-

nite; Put and Lubim were thy helpers.

hiopia fortitudo et Aegyptus et non est finis Africa et Lybies fuerunt in auxilio tuo 3:10 sed et ipsa in transmigrationem 3:10 Yet was she carried away, she went into captivity: her

young children also were dashed in pieces at the top of all the streets: and they cast lots for her honourable men, and

ducta est in captiuitatem paruuli eius elisi sunt in capite om- all her great men were bound in chains. nium uiarum et super inclitos eius miserunt sortem et omnes optimates eius confixi sunt in conpedibus 3:11 et tu ergo 3:11 Thou also shalt be drunken: thou shalt be hid, thou

also shalt seek strength because of the enemy.

inebriaberis eris despecta et tu quaeres auxilium ab inimico 3:12 omnes munitiones tuae sicuti ficus cum grossis suis si 3:12 All thy strong holds shall be like fig trees with the

firstripe figs: if they be shaken, they shall even fall into the mouth of the eater.

concussae fuerint cadent in os comedentis 3:13 ecce popu-

3:13 Behold, thy people in the midst of thee are women: the gates of thy land shall be set wide open unto thine enemies: the fire shall devour thy bars.

lus tuus mulieres in medio tui inimicis tuis adapertione pandentur portae terrae tuae deuorabit ignis uectes tuos 3:14 3:14 Draw thee waters for the siege, fortify thy strong

holds: go into clay, and tread the morter, make strong the brickkiln.

aquam propter obsidionem hauri tibi extrue munitiones tuas intra in lutum et calca subigens tene laterem 3:15 ibi come- 3:15 There shall the fire devour thee; the sword shall cut

thee off, it shall eat thee up like the cankerworm: make thyself many as the cankerworm, make thyself many as

det te ignis peribis gladio deuorabit te ut bruchus congregare the locusts. ut bruchus multiplicare ut lucusta 3:16 plures fecisti nego- 3:16 Thou hast multiplied thy merchants above the stars

of heaven: the cankerworm spoileth, and fleeth away.

tiationes tuas quam stellae sunt caeli bruchus expansus est et auolauit 3:17 custodes tui quasi lucustae et paruuli tui quasi 3:17 Thy crowned are as the locusts, and thy captains as

the great grasshoppers, which camp in the hedges in the cold day, but when the sun ariseth they flee away, and their

lucustae lucustarum quae considunt in sepibus in die frigoris place is not known where they are. sol ortus est et auolauerunt et non est cognitus locus earum ubi fuerint 3:18 dormitauerunt pastores tui rex Assur sepe- 3:18 Thy shepherds slumber, O king of Assyria: thy no-

bles shall dwell in the dust: thy people is scattered upon the mountains, and no man gathereth them.

lientur principes tui latitauit populus tuus in montibus et non est qui congreget 3:19 non est obscura contritio tua pessima 3:19 There is no healing of thy bruise; thy wound is

grievous: all that hear the bruit of thee shall clap the hands over thee: for upon whom hath not thy wickedness passed

est plaga tua omnes qui audierunt auditionem tuam conpres- continually? serunt manum super te quia super quem non transiit malitia

Biblia Sacra Vulgata

tua semper

Prophetia Habacuc

1:1 onus quod uidit Abacuc propheta 1:2 usquequo Do- 1:1 The burden which Habakkuk the prophet did see.

1:2 O LORD, how long shall I cry, and thou wilt not hear! even cry out unto thee of violence, and thou wilt not save!

mine clamabo et non exaudies uociferabor ad te uim patiens et non saluabis 1:3 quare ostendisti mihi iniquitatem 1:3 Why dost thou shew me iniquity, and cause me to be-

hold grievance? for spoiling and violence are before me: and there are that raise up strife and contention.

et laborem uidere praeda et iniustitia contra me et factum est iudicium et contradictio potentior 1:4 propter hoc lace- 1:4 Therefore the law is slacked, and judgment doth never

go forth: for the wicked doth compass about the righteous; therefore wrong judgment proceedeth.

rata est lex et non peruenit usque ad finem iudicium quia impius praeualet aduersus iustum propterea egreditur iudicium peruersum 1:5 aspicite in gentibus et uidete et admi- 1:5 Behold ye among the heathen, and regard, and wonder

marvelously: for I will work a work in your days which ye will not believe, though it be told you.

ramini et obstupescite quia opus factum est in diebus uestris

quod nemo credet cum narrabitur 1:6 quia ecce ego susci- 1:6 For, lo, I raise up the Chaldeans, that bitter and hasty nation, which shall march through the breadth of the land, to possess the dwellingplaces that are not their’s.

tabo Chaldeos gentem amaram et uelocem ambulantem super latitudinem terrae ut possideat tabernacula non sua 1:7 1:7 They are terrible and dreadful: their judgment and

their dignity shall proceed of themselves.

horribilis et terribilis est ex semet ipsa iudicium et onus eius egredietur 1:8 leuiores pardis equi eius et uelociores lupis 1:8 Their horses also are swifter than the leopards, and are

more fierce than the evening wolves: and their horsemen shall spread themselves, and their horsemen shall come

uespertinis et diffundentur equites eius equites namque eius from far; they shall fly as the eagle that hasteth to eat. de longe uenient uolabunt quasi aquila festinans ad come-

dendum 1:9 omnes ad praedam uenient facies eorum uen- 1:9 They shall come all for violence: their faces shall sup up as the east wind, and they shall gather the captivity as the sand.

tus urens et congregabit quasi harenam captiuitatem 1:10

1:10 And they shall scoff at the kings, and the princes shall be a scorn unto them: they shall deride every strong hold; for they shall heap dust, and take it.

et ipse de regibus triumphabit et tyranni ridiculi eius erunt ipse super omnem munitionem ridebit et conportabit aggerem et capiet eam 1:11 tunc mutabitur spiritus et pertransi- 1:11 Then shall his mind change, and he shall pass over,

and offend, imputing this his power unto his god.

bit et corruet haec est fortitudo eius dei sui 1:12 numquid 1:12 Art thou not from everlasting, O LORD my God,

mine Holy One? we shall not die. O LORD, thou hast ordained them for judgment; and, O mighty God, thou hast

non tu a principio Domine Deus meus Sancte meus et non established them for correction. moriemur Domine in iudicium posuisti eum et fortem ut corriperes fundasti eum 1:13 mundi sunt oculi tui ne uideas 1:13 Thou art of purer eyes than to behold evil, and canst

not look on iniquity: wherefore lookest thou upon them that deal treacherously, and holdest thy tongue when the

malum et respicere ad iniquitatem non poteris quare non res- wicked devoureth the man that is more righteous than he? picis super inique agentes et taces deuorante impio iustiorem se 1:14 et facies homines quasi pisces maris et quasi reptile 1:14 And makest men as the fishes of the sea, as the creep-

ing things, that have no ruler over them?

non habens principem 1:15 totum in hamo subleuauit traxit 1:15 They take up all of them with the angle, they catch

them in their net, and gather them in their drag: therefore they rejoice and are glad.

illud in sagena sua et congregauit in rete suo super hoc laetabitur et exultabit 1:16 propterea immolabit sagenae suae 1:16 Therefore they sacrifice unto their net, and burn in-

cense unto their drag; because by them their portion is fat, and their meat plenteous.

et sacrificabit reti suo quia in ipsis incrassata est pars eius et cibus eius electus 1:17 propter hoc ergo expandit sagenam 1:17 Shall they therefore empty their net, and not spare

continually to slay the nations?

suam et semper interficere gentes non parcet

2:1 I will stand upon my watch, and set me upon the tower, 2:1 super custodiam meam stabo et figam gradum super and will watch to see what he will say unto me, and what I shall answer when I am reproved.

munitionem et contemplabor ut uideam quid dicatur mihi et

2:2 And the LORD answered me, and said, Write the vi- quid respondeam ad arguentem me 2:2 et respondit mihi

sion, and make it plain upon tables, that he may run that readeth it.

Dominus et dixit scribe uisum et explana eum super tabulas

2:3 For the vision is yet for an appointed time, but at the ut percurrat qui legerit eum 2:3 quia adhuc uisus procul et

end it shall speak, and not lie: though it tarry, wait for it; because it will surely come, it will not tarry.

apparebit in finem et non mentietur si moram fecerit expecta

2:4 Behold, his soul which is lifted up is not upright in illum quia ueniens ueniet et non tardabit 2:4 ecce qui in-

him: but the just shall live by his faith.

credulus est non erit recta anima eius in semet ipso iustus

2:5 Yea also, because he transgresseth by wine, he is a autem in fide sua uiuet 2:5 et quomodo uinum potantem proud man, neither keepeth at home, who enlargeth his desire as hell, and is as death, and cannot be satisfied, but gathereth unto him all nations, and heapeth unto him all people:

decipit sic erit uir superbus et non decorabitur qui dilatauit quasi infernus animam suam et ipse quasi mors et non adimpletur et congregabit ad se omnes gentes et coacerua-

2:6 Shall not all these take up a parable against him, and bit ad se omnes populos 2:6 numquid non omnes isti super a taunting proverb against him, and say, Woe to him that increaseth that which is not his! how long? and to him

that ladeth himself with thick clay! eum parabolam sument et loquellam enigmatum eius et dicetur uae ei qui multiplicat non sua usquequo et adgrauat

2:7 Shall they not rise up suddenly that shall bite thee, and contra se densum lutum 2:7 numquid non repente consurawake that shall vex thee, and thou shalt be for booties unto them?

gent qui mordeant te et suscitabuntur lacerantes te et eris in

2:8 Because thou hast spoiled many nations, all the rem- rapinam eis 2:8 quia tu spoliasti gentes multas spoliabunt

nant of the people shall spoil thee; because of men’s blood, and for the violence of the land, of the city, and of all that

dwell therein. te omnes qui reliqui fuerint de populis propter sanguinem hominis et iniquitatem terrae ciuitatis et omnium habitan-

2:9 Woe to him that coveteth an evil covetousness to his tium in ea 2:9 uae qui congregat auaritiam malam domui house, that he may set his nest on high, that he may be delivered from the power of evil!

suae ut sit in excelso nidus eius et liberari se putat de manu

2:10 Thou hast consulted shame to thy house by cutting mali 2:10 cogitasti confusionem domui tuae concidisti po-

off many people, and hast sinned against thy soul.

2:11 For the stone shall cry out of the wall, and the beam pulos multos et peccauit anima tua 2:11 quia lapis de pa-

out of the timber shall answer it.

riete clamabit et lignum quod inter iuncturas aedificiorum

2:12 Woe to him that buildeth a town with blood, and sta- est respondebit 2:12 uae qui aedificat ciuitatem in sanguini-

blisheth a city by iniquity!

2:13 Behold, is it not of the LORD of hosts that the people bus et praeparat urbem in iniquitate 2:13 numquid non haec

shall labour in the very fire, and the people shall weary themselves for very vanity?

a Domino sunt exercituum laborabunt enim populi in multo

2:14 For the earth shall be filled with the knowledge of the igni et gentes in uacuum et deficient 2:14 quia replebitur glory of the LORD, as the waters cover the sea.

terra ut cognoscat gloriam Domini quasi aquae operientes

2:15 Woe unto him that giveth his neighbour drink, that mare 2:15 uae qui potum dat amico suo mittens fel suum et puttest thy bottle to him, and makest him drunken also, that thou mayest look on their nakedness!

inebrians ut aspiciat nuditatem eius 2:16 repletus est igno-

2:16 Thou art filled with shame for glory: drink thou also, and let thy foreskin be uncovered: the cup of the LORD’s right hand shall be turned unto thee, and shameful spewing shall be on thy glory.

Prophetia Habacuc

minia pro gloria bibe tu quoque et consopire circumdabit te calix dexterae Domini et uomitus ignominiae super gloriam tuam 2:17 quia iniquitas Libani operiet te et uastitas ani- 2:17 For the violence of Lebanon shall cover thee, and the

spoil of beasts, which made them afraid, because of men’s blood, and for the violence of the land, of the city, and of

malium deterrebit eos de sanguinibus hominis et iniquitate all that dwell therein. terrae et ciuitatis et omnium habitantium in ea 2:18 quid 2:18 What profiteth the graven image that the maker

thereof hath graven it; the molten image, and a teacher of lies, that the maker of his work trusteth therein, to make

prodeet sculptile quia sculpsit illud fictor suus conflatile et dumb idols? imaginem falsam quia sperauit in figmento fictor eius ut faceret simulacra muta 2:19 uae qui dicit ligno expergiscere 2:19 Woe unto him that saith to the wood, Awake; to the

dumb stone, Arise, it shall teach! Behold, it is laid over with gold and silver, and there is no breath at all in the

surge lapidi tacenti numquid ipse docere poterit ecce iste co- midst of it. opertus est auro et argento et omnis spiritus non est in uisceribus eius 2:20 Dominus autem in templo sancto suo sileat 2:20 But the LORD is in his holy temple: let all the earth

keep silence before him.

a facie eius omnis terra

3:1 oratio Abacuc prophetae pro ignorationibus 3:2 Do- 3:1 A prayer of Habakkuk the prophet upon Shigionoth.

3:2 O LORD, I have heard thy speech, and was afraid: O LORD, revive thy work in the midst of the years, in the midst of the years make known; in wrath remember mercy.

mine audiui auditionem tuam et timui Domine opus tuum in medio annorum uiuifica illud in medio annorum notum facies cum iratus fueris misericordiae recordaberis 3:3 Deus 3:3 God came from Teman, and the Holy One from mount

Paran. Selah. His glory covered the heavens, and the earth was full of his praise.

ab austro ueniet et Sanctus de monte Pharan semper operuit caelos gloria eius et laudis eius plena est terra 3:4 splen- 3:4 And his brightness was as the light; he had horns com-

ing out of his hand: and there was the hiding of his power.

dor eius ut lux erit cornua in manibus eius ibi abscondita est fortitudo eius 3:5 ante faciem eius ibit mors et egredie- 3:5 Before him went the pestilence, and burning coals

went forth at his feet.

tur diabolus ante pedes eius 3:6 stetit et mensus est terram 3:6 He stood, and measured the earth: he beheld, and

drove asunder the nations; and the everlasting mountains were scattered, the perpetual hills did bow: his ways are

aspexit et dissoluit gentes et contriti sunt montes saeculi in- everlasting. curuati sunt colles mundi ab itineribus aeternitatis eius 3:7 3:7 I saw the tents of Cushan in affliction: and the curtains

of the land of Midian did tremble.

pro iniquitate uidi tentoria Aethiopiae turbabuntur pelles terrae Madian 3:8 numquid in fluminibus iratus es Domine 3:8 Was the LORD displeased against the rivers? was

thine anger against the rivers? was thy wrath against the sea, that thou didst ride upon thine horses and thy chariots

aut in fluminibus furor tuus uel in mari indignatio tua quia of salvation? ascendes super equos tuos et quadrigae tuae saluatio 3:9 3:9 Thy bow was made quite naked, according to the oaths

of the tribes, even thy word. Selah. Thou didst cleave the earth with rivers.

suscitans suscitabis arcum tuum iuramenta tribubus quae locutus es semper fluuios scindes terrae 3:10 uiderunt te et 3:10 The mountains saw thee, and they trembled: the over-

flowing of the water passed by: the deep uttered his voice, and lifted up his hands on high.

doluerunt montes gurges aquarum transiit dedit abyssus uocem suam altitudo manus suas leuauit 3:11 sol et luna stete- 3:11 The sun and moon stood still in their habitation: at

the light of thine arrows they went, and at the shining of thy glittering spear.

runt in habitaculo suo in luce sagittarum tuarum ibunt in splendore fulgurantis hastae tuae 3:12 in fremitu conculcabis 3:12 Thou didst march through the land in indignation,

thou didst thresh the heathen in anger.

Biblia Sacra Vulgata

3:13 Thou wentest forth for the salvation of thy people, terram in furore obstupefacies gentes 3:13 egressus es in

even for salvation with thine anointed; thou woundedst the head out of the house of the wicked, by discovering the

foundation unto the neck. Selah. salutem populi tui in salutem cum christo tuo percussisti ca-

put de domo impii denudasti fundamentum usque ad collum

3:14 Thou didst strike through with his staves the head of semper 3:14 maledixisti sceptris eius capiti bellatorum eius

his villages: they came out as a whirlwind to scatter me: their rejoicing was as to devour the poor secretly.

uenientibus ut turbo ad dispergendum me exultatio eorum

3:15 Thou didst walk through the sea with thine horses, sicut eius qui deuorat pauperem in abscondito 3:15 uiam

through the heap of great waters.

3:16 When I heard, my belly trembled; my lips quivered fecisti in mari equis tuis in luto aquarum multarum 3:16

at the voice: rottenness entered into my bones, and I trembled in myself, that I might rest in the day of trouble: when he cometh up unto the people, he will invade them with his troops.

audiui et conturbatus est uenter meus ad uocem contremuerunt labia mea ingrediatur putredo in ossibus meis et subter me scateat ut requiescam in die tribulationis ut ascendam ad

3:17 Although the fig tree shall not blossom, neither shall populum accinctum nostrum 3:17 ficus enim non florebit et fruit be in the vines; the labour of the olive shall fail, and the fields shall yield no meat; the flock shall be cut off

from the fold, and there shall be no herd in the stalls: non erit germen in uineis mentietur opus oliuae et arua non

adferent cibum abscidetur de ouili pecus et non erit armen-

3:18 Yet I will rejoice in the LORD, I will joy in the God tum in praesepibus 3:18 ego autem in Domino gaudebo ex-

of my salvation.

3:19 The LORD God is my strength, and he will make my ultabo in Deo Iesu meo 3:19 Dominus Deus fortitudo mea

feet like hinds’ feet, and he will make me to walk upon mine high places. To the chief singer on my stringed in-

struments. et ponet pedes meos quasi ceruorum et super excelsa mea deducet me uictori in psalmis canentem

Prophetia Sophoniae

1:1 uerbum Domini quod factum est ad Sofoniam filium 1:1 The word of the LORD which came unto Zephaniah

the son of Cushi, the son of Gedaliah, the son of Amariah, the son of Hizkiah, in the days of Josiah the son of Amon,

Chusi filium Godoliae filii Amariae filii Ezechiae in die- king of Judah.

bus Iosiae filii Amon regis Iuda 1:2 congregans congre- 1:2 I will utterly consume all things from off the land, saith

the LORD.

gabo omnia a facie terrae dicit Dominus 1:3 congregans 1:3 I will consume man and beast; I will consume the

fowls of the heaven, and the fishes of the sea, and the stumbling blocks with the wicked: and I will cut off man from

hominem et pecus congregans uolatile caeli et pisces ma- off the land, saith the LORD. ris et ruinae impiorum erunt et disperdam homines a facie terrae dicit Dominus 1:4 et extendam manum meam super 1:4 I will also stretch out mine hand upon Judah, and

upon all the inhabitants of Jerusalem; and I will cut off the remnant of Baal from this place, and the name of the

Iudam et super omnes habitantes Hierusalem et disperdam Chemarims with the priests; de loco hoc reliquias Baal et nomina aedituorum cum sacerdotibus 1:5 et eos qui adorant super tecta militiam caeli 1:5 And them that worship the host of heaven upon the

housetops; and them that worship and that swear by the LORD, and that swear by Malcham;

et adorant et iurant in Domino et iurant in Melchom 1:6 et

1:6 And them that are turned back from the LORD; and those that have not sought the LORD, nor enquired for him.

qui auertuntur de post tergum Domini et qui non quaesierunt Dominum nec inuestigauerunt eum 1:7 silete a facie 1:7 Hold thy peace at the presence of the Lord GOD: for

the day of the LORD is at hand: for the LORD hath prepared a sacrifice, he hath bid his guests.

Domini Dei quia iuxta est dies Domini quia praeparauit Dominus hostiam sanctificauit uocatos suos 1:8 et erit in die 1:8 And it shall come to pass in the day of the LORD’s

sacrifice, that I will punish the princes, and the king’s children, and all such as are clothed with strange apparel.

hostiae Domini uisitabo super principes et super filios regis et super omnes qui induti sunt ueste peregrina 1:9 et uisi- 1:9 In the same day also will I punish all those that leap

on the threshold, which fill their masters’ houses with violence and deceit.

tabo omnem qui arroganter ingreditur super limen in die illa qui conplent domum Domini Dei sui iniquitate et dolo 1:10 1:10 And it shall come to pass in that day, saith the LORD,

that there shall be the noise of a cry from the fish gate, and an howling from the second, and a great crashing from the

et erit in die illa dicit Dominus uox clamoris a porta Piscium hills. et ululatus a secunda et contritio magna a collibus 1:11 ulu- 1:11 Howl, ye inhabitants of Maktesh, for all the merchant

people are cut down; all they that bear silver are cut off.

late habitatores pilae conticuit omnis populus Chanaan disperierunt omnes inuoluti argento 1:12 et erit in tempore illo 1:12 And it shall come to pass at that time, that I will

search Jerusalem with candles, and punish the men that are settled on their lees: that say in their heart, The LORD

scrutabor Hierusalem in lucernis et uisitabo super uiros de- will not do good, neither will he do evil. fixos in fecibus suis qui dicunt in cordibus suis non faciet

bene Dominus et non faciet male 1:13 et erit fortitudo eo- 1:13 Therefore their goods shall become a booty, and their houses a desolation: they shall also build houses, but not inhabit them; and they shall plant vineyards, but not drink

rum in direptionem et domus eorum in desertum et aedifica- the wine thereof. bunt domos et non habitabunt et plantabunt uineas et non bibent uinum earum 1:14 iuxta est dies Domini magnus iuxta 1:14 The great day of the LORD is near, it is near, and

hasteth greatly, even the voice of the day of the LORD: the mighty man shall cry there bitterly.

et uelox nimis uox diei Domini amara tribulabitur ibi fortis 1:15 dies irae dies illa dies tribulationis et angustiae dies cala- 1:15 That day is a day of wrath, a day of trouble and dis-

tress, a day of wasteness and desolation, a day of darkness and gloominess, a day of clouds and thick darkness,

mitatis et miseriae dies tenebrarum et caliginis dies nebulae et turbinis 1:16 dies tubae et clangoris super ciuitates mu- 1:16 A day of the trumpet and alarm against the fenced

cities, and against the high towers.

Biblia Sacra Vulgata

1:17 And I will bring distress upon men, that they shall nitas et super angulos excelsos 1:17 et tribulabo homines et

walk like blind men, because they have sinned against the LORD: and their blood shall be poured out as dust, and

their flesh as the dung. ambulabunt ut caeci quia Domino peccauerunt et effundetur

sanguis eorum sicut humus et corpus eorum sicut stercora

1:18 Neither their silver nor their gold shall be able to de- 1:18 sed et argentum eorum et aurum eorum non poterit liliver them in the day of the LORD’s wrath; but the whole land shall be devoured by the fire of his jealousy: for he shall make even a speedy riddance of all them that dwell in the land.

berare eos in die irae Domini in igne zeli eius deuorabitur omnis terra quia consummationem cum festinatione faciet cunctis habitantibus terram

2:1 Gather yourselves together, yea, gather together, O na- 2:1 conuenite congregamini gens non amabilis 2:2 priustion not desired; 2:2 Before the decree bring forth, before the day pass as the chaff, before the fierce anger of the LORD come upon you, before the day of the LORD’s anger come upon you.

quam pariat iussio quasi puluerem transeuntem diem antequam ueniat super uos ira furoris Domini antequam ueniat

2:3 Seek ye the LORD, all ye meek of the earth, which super uos dies furoris Domini 2:3 quaerite Dominum om-

have wrought his judgment; seek righteousness, seek meekness: it may be ye shall be hid in the day of the

LORD’s anger. nes mansueti terrae qui iudicium eius estis operati quaerite

iustum quaerite mansuetum si quo modo abscondamini in

2:4 For Gaza shall be forsaken, and Ashkelon a desolation: die furoris Domini 2:4 quia Gaza destructa erit et Ascalon they shall drive out Ashdod at the noon day, and Ekron shall be rooted up.

in desertum Azotum in meridie eicient et Accaron eradicabi-

2:5 Woe unto the inhabitants of the sea coast, the nation tur 2:5 uae qui habitatis funiculum maris gens perditorum

of the Cherethites! the word of the LORD is against you; O Canaan, the land of the Philistines, I will even destroy

thee, that there shall be no inhabitant. uerbum Domini super uos Chanaan terra Philisthinorum et 2:6 And the sea coast shall be dwellings and cottages for disperdam te ita ut non sit inhabitator 2:6 et erit funiculus

shepherds, and folds for flocks.

2:7 And the coast shall be for the remnant of the house maris requies pastorum et caulae pecorum 2:7 et erit fu-

of Judah; they shall feed thereupon: in the houses of Ashkelon shall they lie down in the evening: for the LORD their God shall visit them, and turn away their captivity. niculus eius qui remanserit de domo Iuda ibi pascentur in

domibus Ascalonis ad uesperam requiescent quia uisitabit

2:8 I have heard the reproach of Moab, and the revilings eos Dominus Deus eorum et auertet captiuitatem eorum 2:8

of the children of Ammon, whereby they have reproached my people, and magnified themselves against their border.

audiui obprobrium Moab et blasphemias filiorum Ammon quae exprobrauerunt populo meo et magnificati sunt super

2:9 Therefore as I live, saith the LORD of hosts, the God terminos eorum 2:9 propterea uiuo ego dicit Dominus ex-

of Israel, Surely Moab shall be as Sodom, and the children of Ammon as Gomorrah, even the breeding of nettles, and saltpits, and a perpetual desolation: the residue of my people shall spoil them, and the remnant of my people shall possess them.

ercituum Deus Israhel quia Moab ut Sodoma erit et filii Ammon quasi Gomorra siccitas spinarum et acerui salis et desertum usque in aeternum reliquiae populi mei diripient illos

2:10 This shall they have for their pride, because they have residui gentis meae possidebunt eos 2:10 hoc eis eueniet reproached and magnified themselves against the people of the LORD of hosts.

pro superbia sua quia blasphemauerunt et magnificati sunt

2:11 The LORD will be terrible unto them: for he will super populum Domini exercituum 2:11 horribilis Domi-

famish all the gods of the earth; and men shall worship him, every one from his place, even all the isles of the

heathen. nus super eos et adtenuabit omnes deos terrae et adorabunt 2:12 Ye Ethiopians also, ye shall be slain by my sword. eum uir de loco suo omnes insulae gentium 2:12 sed et uos

Prophetia Sophoniae

Aethiopes interfecti gladio meo eritis 2:13 et extendet ma- 2:13 And he will stretch out his hand against the north,

and destroy Assyria; and will make Nineveh a desolation, and dry like a wilderness.

num suam super aquilonem et perdet Assur et ponet speciosam in solitudinem et in inuium et quasi desertum 2:14 et 2:14 And flocks shall lie down in the midst of her, all the

beasts of the nations: both the cormorant and the bittern shall lodge in the upper lintels of it; their voice shall sing in the windows; desolation shall be in the thresholds; for he shall uncover the cedar work.

accubabunt in medio eius greges omnes bestiae gentium et onocrotalus et ericius in liminibus eius morabuntur uox cantantis in fenestra coruus in superliminari quoniam adtenuabo robur eius 2:15 haec est ciuitas gloriosa habitans in confi- 2:15 This is the rejoicing city that dwelt carelessly, that

said in her heart, I am, and there is none beside me: how is she become a desolation, a place for beasts to lie down in! every one that passeth by her shall hiss, and wag his hand.

dentia quae dicebat in corde suo ego sum et extra me non est alia amplius quomodo facta est in desertum cubile bestiae omnis qui transit per eam sibilabit et mouebit manum suam

3:1 uae prouocatrix et redeepta ciuitas columba 3:2 non 3:1 Woe to her that is filthy and polluted, to the oppressing city! 3:2 She obeyed not the voice; she received not correction; she trusted not in the LORD; she drew not near to her God.

audiuit uocem et non suscepit disciplinam in Domino non

est confisa ad Deum suum non adpropiauit 3:3 principes 3:3 Her princes within her are roaring lions; her judges are

evening wolves; they gnaw not the bones till the morrow.

eius in medio eius quasi leones rugientes iudices eius lupi uespere non relinquebant in mane 3:4 prophetae eius ue- 3:4 Her prophets are light and treacherous persons: her

priests have polluted the sanctuary, they have done violence to the law.

sani uiri infideles sacerdotes eius polluerunt sanctum iniuste egerunt contra legem 3:5 Dominus iustus in medio eius non 3:5 The just LORD is in the midst thereof; he will not

do iniquity: every morning doth he bring his judgment to light, he faileth not; but the unjust knoweth no shame.

faciet iniquitatem mane mane iudicium suum dabit in luce et non abscondetur nesciuit autem iniquus confusionem 3:6 3:6 I have cut off the nations: their towers are desolate; I

made their streets waste, that none passeth by: their cities are destroyed, so that there is no man, that there is none

disperdi gentes et dissipati sunt anguli earum desertas feci inhabitant. uias eorum dum non est qui transeat desolatae sunt ciuitates eorum non remanente uiro nec ullo habitatore 3:7 dixi at- 3:7 I said, Surely thou wilt fear me, thou wilt receive in-

struction; so their dwelling should not be cut off, howsoever I punished them: but they rose early, and corrupted

tamen timebis me suscipies disciplinam et non peribit habi- all their doings. taculum eius propter omnia in quibus uisitaui eam uerumtamen diluculo surgentes corruperunt omnes cogitationes suas 3:8 quapropter expecta me dicit Dominus in die resurrec- 3:8 Therefore wait ye upon me, saith the LORD, until the

day that I rise up to the prey: for my determination is to gather the nations, that I may assemble the kingdoms, to pour upon them mine indignation, even all my fierce anger: for all the earth shall be devoured with the fire of my jealousy.

tionis meae in futurum quia iudicium meum ut congregem gentes et colligam regna ut effundam super eas indignationem meam omnem iram furoris mei in igne enim zeli mei deuorabitur omnis terra 3:9 quia tunc reddam populis la- 3:9 For then will I turn to the people a pure language, that

they may all call upon the name of the LORD, to serve him with one consent.

bium electum ut uocent omnes in nomine Domini et seruiant ei umero uno 3:10 ultra flumina Aethiopiae inde supplices 3:10 From beyond the rivers of Ethiopia my suppliants,

even the daughter of my dispersed, shall bring mine offering.

mei filii dispersorum meorum deferent munus mihi 3:11 in

3:11 In that day shalt thou not be ashamed for all thy doings, wherein thou hast transgressed against me: for then I will take away out of the midst of thee them that rejoice in thy pride, and thou shalt no more be haughty because of my holy mountain.

MCCCXV

Biblia Sacra Vulgata

die illa non confunderis super cunctis adinuentionibus tuis quibus praeuaricata es in me quia tunc auferam de medio tui magniloquos superbiae tuae et non adicies exaltari am-

3:12 I will also leave in the midst of thee an afflicted and plius in monte sancto meo 3:12 et derelinquam in medio poor people, and they shall trust in the name of the LORD.

tui populum pauperem et egenum et sperabunt in nomine

3:13 The remnant of Israel shall not do iniquity, nor speak Domini 3:13 reliquiae Israhel non facient iniquitatem nec

lies; neither shall a deceitful tongue be found in their mouth: for they shall feed and lie down, and none shall

make them afraid. loquentur mendacium et non inuenietur in ore eorum lin-

gua dolosa quoniam ipsi pascentur et accubabunt et non erit

3:14 Sing, O daughter of Zion; shout, O Israel; be glad qui exterreat 3:14 lauda filia Sion iubilate Israhel laetare et

and rejoice with all the heart, O daughter of Jerusalem.

3:15 The LORD hath taken away thy judgments, he hath exulta in omni corde filia Hierusalem 3:15 abstulit Domi-

cast out thine enemy: the king of Israel, even the LORD, is in the midst of thee: thou shalt not see evil any more.

nus iudicium tuum auertit inimicos tuos rex Israhel Domi-

3:16 In that day it shall be said to Jerusalem, Fear thou nus in medio tui non timebis malum ultra 3:16 in die illa not: and to Zion, Let not thine hands be slack.

dicetur Hierusalem noli timere Sion non dissoluantur ma-

3:17 The LORD thy God in the midst of thee is mighty; he nus tuae 3:17 Dominus Deus tuus in medio tui Fortis ipse

will save, he will rejoice over thee with joy; he will rest in his love, he will joy over thee with singing.

saluabit gaudebit super te in laetitia silebit in dilectione tua

3:18 I will gather them that are sorrowful for the solemn exultabit super te in laude 3:18 nugas qui a lege recesserant

assembly, who are of thee, to whom the reproach of it was a burden.

congregabo quia ex te erant ut non ultra habeas super eis

3:19 Behold, at that time I will undo all that afflict thee: obprobrium 3:19 ecce ego interficiam omnes qui adflixeand I will save her that halteth, and gather her that was driven out; and I will get them praise and fame in every

land where they have been put to shame. runt te in tempore illo et saluabo claudicantem et eam quae

eiecta fuerat congregabo et ponam eos in laudem et in no-

3:20 At that time will I bring you again, even in the time men in omni terra confusionis eorum 3:20 in tempore illo

that I gather you: for I will make you a name and a praise among all people of the earth, when I turn back your cap-

tivity before your eyes, saith the LORD. quo adducam uos et in tempore quo congregabo uos dabo enim uos in nomen et in laudem omnibus populis terrae cum conuertero captiuitatem uestram coram oculis uestris dicit Dominus

Prophetia Aggaei

1:1 in anno secundo Darii regis in mense sexto in die una 1:1 In the second year of Darius the king, in the sixth

month, in the first day of the month, came the word of the LORD by Haggai the prophet unto Zerubbabel the son of Shealtiel, governor of Judah, and to Joshua the son of Josedech, the high priest, saying,

mensis factum est uerbum Domini in manu Aggei prophetae ad Zorobabel filium Salathihel ducem Iuda et ad Iesum

filium Iosedech sacerdotem magnum dicens 1:2 haec ait 1:2 Thus speaketh the LORD of hosts, saying, This people say, The time is not come, the time that the LORD’s house should be built.

Dominus exercituum dicens populus iste dicit nondum uenit tempus domus Domini aedificandae 1:3 et factum est 1:3 Then came the word of the LORD by Haggai the

prophet, saying,

uerbum Domini in manu Aggei prophetae dicens 1:4 num- 1:4 Is it time for you, O ye, to dwell in your cieled houses,

and this house lie waste?

quid tempus uobis est ut habitetis in domibus laqueatis et domus ista deserta 1:5 et nunc haec dicit Dominus exercituum 1:5 Now therefore thus saith the LORD of hosts; Consider

your ways.

ponite corda uestra super uias uestras 1:6 seminastis mul- 1:6 Ye have sown much, and bring in little; ye eat, but ye

have not enough; ye drink, but ye are not filled with drink; ye clothe you, but there is none warm; and he that earneth

tum et intulistis parum comedistis et non estis satiati bibistis wages earneth wages to put it into a bag with holes. et non estis inebriati operuistis uos et non estis calefacti et qui mercedes congregauit misit eas in sacculum pertusum 1:7 haec dicit Dominus exercituum ponite corda uestra su- 1:7 Thus saith the LORD of hosts; Consider your ways. per uias uestras 1:8 ascendite in montem portate lignum et 1:8 Go up to the mountain, and bring wood, and build the

house; and I will take pleasure in it, and I will be glorified, saith the LORD.

aedificate domum et acceptabilis mihi erit et glorificabor dicit Dominus 1:9 respexistis ad amplius et ecce factum est 1:9 Ye looked for much, and, lo it came to little; and when

ye brought it home, I did blow upon it. Why? saith the LORD of hosts. Because of mine house that is waste, and

minus et intulistis in domum et exsuflaui illud quam ob cau- ye run every man unto his own house. sam dicit Dominus exercituum quia domus mea deserta est et uos festinatis unusquisque in domum suam 1:10 propter 1:10 Therefore the heaven over you is stayed from dew,

and the earth is stayed from her fruit.

hoc super uos prohibiti sunt caeli ne darent rorem et terra prohibita est ne daret germen suum 1:11 et uocaui siccita- 1:11 And I called for a drought upon the land, and upon

the mountains, and upon the corn, and upon the new wine, and upon the oil, and upon that which the ground bringeth forth, and upon men, and upon cattle, and upon all the labour of the hands.

tem super terram et super montes et super triticum et super uinum et super oleum et quaecumque profert humus et super homines et super iumenta et super omnem laborem manuum 1:12 et audiuit Zorobabel filius Salathihel et Iesus filius Iose- 1:12 Then Zerubbabel the son of Shealtiel, and Joshua the

son of Josedech, the high priest, with all the remnant of the people, obeyed the voice of the LORD their God, and the words of Haggai the prophet, as the LORD their God had sent him, and the people did fear before the LORD.

dech sacerdos magnus et omnes reliquiae populi uocem Dei sui et uerba Aggei prophetae sicut misit eum Dominus Deus eorum ad ipsos et timuit populus a facie Domini 1:13 et 1:13 Then spake Haggai the LORD’s messenger in the

LORD’s message unto the people, saying, I am with you, saith the LORD.

dixit Aggeus nuntius Domini de nuntiis Domini populo dicens ego uobiscum dicit Dominus 1:14 et suscitauit Domi- 1:14 And the LORD stirred up the spirit of Zerubbabel

the son of Shealtiel, governor of Judah, and the spirit of Joshua the son of Josedech, the high priest, and the spirit of all the remnant of the people; and they came and did work in the house of the LORD of hosts, their God,

nus spiritum Zorobabel filii Salathihel ducis Iuda et spiritum Iesu filii Iosedech sacerdotis magni et spiritum reliquorum

Biblia Sacra Vulgata

de omni populo et ingressi sunt et faciebant opus in domo Domini exercituum Dei sui

2:1 In the four and twentieth day of the sixth month, in the 2:1 in die uicesima et quarta mensis in sexto mense in second year of Darius the king. 2:2 In the seventh month, in the one and twentieth day of anno secundo Darii regis 2:2 in septimo mense uicesima the month, came the word of the LORD by the prophet Haggai, saying,

et prima mensis factum est uerbum Domini in manu Ag-

2:3 Speak now to Zerubbabel the son of Shealtiel, gover- gei prophetae dicens 2:3 loquere ad Zorobabel filium Sanor of Judah, and to Joshua the son of Josedech, the high priest, and to the residue of the people, saying,

lathihel ducem Iuda et ad Iesum filium Iosedech sacerdotem

2:4 Who is left among you that saw this house in her first magnum et ad reliquos populi dicens 2:4 quis in uobis est

glory? and how do ye see it now? is it not in your eyes in comparison of it as nothing?

derelictus qui uidit domum istam in gloria sua prima et quid uos uidetis hanc nunc numquid non ita est quasi non sit in

2:5 Yet now be strong, O Zerubbabel, saith the LORD; and oculis uestris 2:5 et nunc confortare Zorobabel dicit Domi-

be strong, O Joshua, son of Josedech, the high priest; and be strong, all ye people of the land, saith the LORD, and

work: for I am with you, saith the LORD of hosts: nus et confortare Iesu fili Iosedech sacerdos magne et confortare omnis popule terrae dicit Dominus exercituum et facite quoniam ego uobiscum sum dicit Dominus exercituum

2:6 According to the word that I covenanted with you 2:6 uerbum quod placui uobiscum cum egrederemini de terra

when ye came out of Egypt, so my spirit remaineth among you: fear ye not.

Aegypti et spiritus meus erit in medio uestrum nolite timere

2:7 For thus saith the LORD of hosts; Yet once, it is a little 2:7 quia haec dicit Dominus exercituum adhuc unum mowhile, and I will shake the heavens, and the earth, and the sea, and the dry land;

dicum est et ego commouebo caelum et terram et mare et

2:8 And I will shake all nations, and the desire of all na- aridam 2:8 et mouebo omnes gentes et ueniet desideratus tions shall come: and I will fill this house with glory, saith the LORD of hosts.

cunctis gentibus et implebo domum istam gloria dicit Domi-

2:9 The silver is mine, and the gold is mine, saith the nus exercituum 2:9 meum est argentum et meum est aurum

LORD of hosts. 2:10 The glory of this latter house shall be greater than of dicit Dominus exercituum 2:10 magna erit gloria domus isthe former, saith the LORD of hosts: and in this place will I give peace, saith the LORD of hosts.

tius nouissimae plus quam primae dicit Dominus exercituum

2:11 In the four and twentieth day of the ninth month, in et in loco isto dabo pacem dicit Dominus exercituum 2:11

the second year of Darius, came the word of the LORD by Haggai the prophet, saying,

in uicesima et quarta noni mensis in anno secundo Darii fac-

2:12 Thus saith the LORD of hosts; Ask now the priests tum est uerbum Domini ad Aggeum prophetam dicens 2:12

concerning the law, saying,

haec dicit Dominus exercituum interroga sacerdotes legem

2:13 If one bear holy flesh in the skirt of his garment, and dicens 2:13 si tulerit homo carnem sanctificatam in ora ue-

with his skirt do touch bread, or pottage, or wine, or oil, or any meat, shall it be holy? And the priests answered and

said, No. stimenti sui et tetigerit de summitate eius panem aut pulmen-

tum aut uinum aut oleum aut omnem cibum numquid sancti-

2:14 Then said Haggai, If one that is unclean by a dead ficabitur respondentes autem sacerdotes dixerunt non 2:14

body touch any of these, shall it be unclean? And the priests answered and said, It shall be unclean.

et dixit Aggeus si tetigerit pollutus in anima ex omnibus his numquid contaminabitur et responderunt sacerdotes et dixe-

Prophetia Aggaei

runt contaminabitur 2:15 et respondit Aggeus et dixit sic 2:15 Then answered Haggai, and said, So is this people,

and so is this nation before me, saith the LORD; and so is every work of their hands; and that which they offer there

populus iste et sic gens ista ante faciem meam dicit Domi- is unclean. nus et sic omne opus manuum eorum et omnia quae obtulerint ibi contaminata erunt 2:16 et nunc ponite corda uestra 2:16 And now, I pray you, consider from this day and up-

ward, from before a stone was laid upon a stone in the temple of the LORD:

a die hac et supra antequam poneretur lapis super lapidem in templo Domini 2:17 cum accederetis ad aceruum uiginti 2:17 Since those days were, when one came to an heap of

twenty measures, there were but ten: when one came to the pressfat for to draw out fifty vessels out of the press,

modiorum et fierent decem intraretis ad torcular ut exprime- there were but twenty. retis quinquaginta lagoenas et fiebant uiginti 2:18 percussi 2:18 I smote you with blasting and with mildew and with

hail in all the labours of your hands; yet ye turned not to me, saith the LORD.

uos uento urente et aurugine et grandine omnia opera manuum uestrarum et non fuit in uobis qui reuerteretur ad me dicit Dominus 2:19 ponite corda uestra ex die ista et in fu- 2:19 Consider now from this day and upward, from the

four and twentieth day of the ninth month, even from the day that the foundation of the LORD’s temple was laid,

turum a die uicesima et quarta noni mensis a die qua fun- consider it. damenta iacta sunt templi Domini ponite super cor uestrum 2:20 numquid iam semen in germine est et adhuc uinea et fi- 2:20 Is the seed yet in the barn? yea, as yet the vine, and

the fig tree, and the pomegranate, and the olive tree, hath not brought forth: from this day will I bless you.

cus et malogranatum et lignum oliuae non floruit ex die ista benedicam 2:21 et factum est uerbum Domini secundo ad 2:21 And again the word of the LORD came unto Haggai

in the four and twentieth day of the month, saying,

Aggeum in uicesima et quarta mensis dicens 2:22 loquere 2:22 Speak to Zerubbabel, governor of Judah, saying, I

will shake the heavens and the earth;

ad Zorobabel ducem Iuda dicens ego mouebo caelum pariter et terram 2:23 et subuertam solium regnorum et conteram 2:23 And I will overthrow the throne of kingdoms, and I

will destroy the strength of the kingdoms of the heathen; and I will overthrow the chariots, and those that ride in them; and the horses and their riders shall come down, every one by the sword of his brother.

fortitudinem regni gentium et subuertam quadrigam et ascensorem eius et descendent equi et ascensores eorum uir in gladio fratris sui 2:24 in die illo dicit Dominus exerci- 2:24 In that day, saith the LORD of hosts, will I take thee,

O Zerubbabel, my servant, the son of Shealtiel, saith the LORD, and will make thee as a signet: for I have chosen

tuum adsumam te Zorobabel fili Salathihel serue meus dicit thee, saith the LORD of hosts. Dominus et ponam te quasi signaculum quia te elegi dicit Dominus exercituum

Biblia Sacra Vulgata

MCCCXX

Prophetia Zachariae

1:1 in mense octauo in anno secundo Darii factum est uer- 1:1 In the eighth month, in the second year of Darius,

came the word of the LORD unto Zechariah, the son of Berechiah, the son of Iddo the prophet, saying,

bum Domini ad Zacchariam filium Barachiae filium Addo prophetam dicens 1:2 iratus est Dominus super patres ue- 1:2 The LORD hath been sore displeased with your fa-

thers.

stros iracundia 1:3 et dices ad eos haec dicit Dominus exer- 1:3 Therefore say thou unto them, Thus saith the LORD

of hosts; Turn ye unto me, saith the LORD of hosts, and I will turn unto you, saith the LORD of hosts.

cituum conuertimini ad me ait Dominus exercituum et conuertar ad uos dicit Dominus exercituum 1:4 ne sitis sicut 1:4 Be ye not as your fathers, unto whom the former

prophets have cried, saying, Thus saith the LORD of hosts; Turn ye now from your evil ways, and from your evil doings: but they did not hear, nor hearken unto me, saith the LORD.

patres uestri ad quos clamabant prophetae priores dicentes haec dicit Dominus exercituum conuertimini de uiis uestris malis et cogitationibus uestris pessimis et non audierunt neque adtenderunt ad me dicit Dominus 1:5 patres uestri ubi 1:5 Your fathers, where are they? and the prophets, do

they live for ever?

sunt et prophetae numquid in sempiternum uiuent 1:6 ue- 1:6 But my words and my statutes, which I commanded

my servants the prophets, did they not take hold of your fathers? and they returned and said, Like as the LORD of hosts thought to do unto us, according to our ways, and according to our doings, so hath he dealt with us.

rumtamen uerba mea et legitima mea quae mandaui seruis meis prophetis numquid non conprehenderunt patres uestros et conuersi sunt et dixerunt sicut cogitauit Dominus exercituum facere nobis secundum uias nostras et secundum adinuentiones nostras fecit nobis 1:7 in die uicesima et quarta 1:7 Upon the four and twentieth day of the eleventh month,

which is the month Sebat, in the second year of Darius, came the word of the LORD unto Zechariah, the son of

undecimo mense sabath in anno secundo Darii factum est Berechiah, the son of Iddo the prophet, saying, uerbum Domini ad Zacchariam filium Barachiae filium Addo prophetam dicens 1:8 uidi per noctem et ecce uir ascen- 1:8 I saw by night, and behold a man riding upon a red

horse, and he stood among the myrtle trees that were in the bottom; and behind him were there red horses, speckled,

dens super equum rufum et ipse stabat inter myrteta quae and white. erant in profundo et post eum equi rufi uarii et albi 1:9 et 1:9 Then said I, O my lord, what are these? And the angel

that talked with me said unto me, I will shew thee what these be.

dixi quid sunt isti domine mi et dixit ad me angelus qui loquebatur in me ego ostendam tibi quid sint haec 1:10 et res- 1:10 And the man that stood among the myrtle trees an-

swered and said, These are they whom the LORD hath sent to walk to and fro through the earth.

pondit uir qui stabat inter myrteta et dixit isti sunt quos misit Dominus ut perambularent terram 1:11 et responderunt an- 1:11 And they answered the angel of the LORD that stood

among the myrtle trees, and said, We have walked to and fro through the earth, and, behold, all the earth sitteth still,

gelo Domini qui stabat inter myrteta et dixerunt perambula- and is at rest. uimus terram et ecce omnis terra habitatur et quiescit 1:12 1:12 Then the angel of the LORD answered and said, O

LORD of hosts, how long wilt thou not have mercy on Jerusalem and on the cities of Judah, against which thou

et respondit angelus Domini et dixit Domine exercituum us- hast had indignation these threescore and ten years? quequo tu non misereberis Hierusalem et urbium Iuda quibus iratus es iste septuagesimus annus est 1:13 et respon- 1:13 And the LORD answered the angel that talked with

me with good words and comfortable words.

dit Dominus angelo qui loquebatur in me uerba bona uerba consolatoria 1:14 et dixit ad me angelus qui loquebatur in 1:14 So the angel that communed with me said unto me,

Cry thou, saying, Thus saith the LORD of hosts; I am jealous for Jerusalem and for Zion with a great jealousy.

me clama dicens haec dicit Dominus exercituum zelatus sum

Biblia Sacra Vulgata

1:15 And I am very sore displeased with the heathen that Hierusalem et Sion zelo magno 1:15 et ira magna ego iraare at ease: for I was but a little displeased, and they helped forward the affliction.

scor super gentes opulentas quia ego iratus sum parum ipsi

1:16 Therefore thus saith the LORD; I am returned to uero adiuuerunt in malum 1:16 propterea haec dicit Domi-

Jerusalem with mercies: my house shall be built in it, saith the LORD of hosts, and a line shall be stretched forth upon

Jerusalem. nus reuertar ad Hierusalem in misericordiis domus mea ae-

dificabitur in ea ait Dominus exercituum et perpendiculum

1:17 Cry yet, saying, Thus saith the LORD of hosts; My extendetur super Hierusalem 1:17 adhuc clama dicens haec

cities through prosperity shall yet be spread abroad; and the LORD shall yet comfort Zion, and shall yet choose

Jerusalem. dicit Dominus exercituum adhuc affluent ciuitates meae bo-

nis et consolabitur Dominus adhuc Sion et eliget adhuc Hie-

1:18 Then lifted I up mine eyes, and saw, and behold four rusalem 1:18 et leuaui oculos meos et uidi et ecce quattuor

horns.

1:19 And I said unto the angel that talked with me, What cornua 1:19 et dixi ad angelum qui loquebatur in me quid

be these? And he answered me, These are the horns which have scattered Judah, Israel, and Jerusalem.

sunt haec et dixit ad me haec sunt cornua quae uentilauerunt

1:20 And the LORD shewed me four carpenters. Iudam et Israhel et Hierusalem 1:20 et ostendit mihi Do-1:21 Then said I, What come these to do? And he spake, minus quattuor fabros 1:21 et dixi quid isti ueniunt facere

saying, These are the horns which have scattered Judah, so that no man did lift up his head: but these are come to fray them, to cast out the horns of the Gentiles, which lifted up their horn over the land of Judah to scatter it.

qui ait dicens haec sunt cornua quae uentilauerunt Iudam per singulos uiros et nemo eorum leuauit caput suum et uenerunt isti deterrere ea ut deiciant cornua gentium quae leuauerunt cornu super terram Iuda ut dispergerent eam

2:1 I lifted up mine eyes again, and looked, and behold a 2:1 et leuaui oculos meos et uidi et ecce uir et in manu man with a measuring line in his hand.

2:2 Then said I, Whither goest thou? And he said unto me, eius funiculus mensorum 2:2 et dixi quo tu uadis et dixit

To measure Jerusalem, to see what is the breadth thereof, and what is the length thereof.

ad me ut metiar Hierusalem et uideam quanta sit latitudo

2:3 And, behold, the angel that talked with me went forth, eius et quanta longitudo eius 2:3 et ecce angelus qui loand another angel went out to meet him,

quebatur in me egrediebatur et angelus alius egrediebatur in

2:4 And said unto him, Run, speak to this young man, say- occursum eius 2:4 et dixit ad eum curre loquere ad puerum

ing, Jerusalem shall be inhabited as towns without walls for the multitude of men and cattle therein:

istum dicens absque muro habitabitur Hierusalem prae mul-

2:5 For I, saith the LORD, will be unto her a wall of fire titudine hominum et iumentorum in medio eius 2:5 et ego

round about, and will be the glory in the midst of her.

ero ei ait Dominus murus ignis in circuitu et in gloria ero

2:6 Ho, ho, come forth, and flee from the land of the north, in medio eius 2:6 o o fugite de terra aquilonis dicit Domi-

saith the LORD: for I have spread you abroad as the four winds of the heaven, saith the LORD.

nus quoniam in quattuor uentos caeli dispersi uos dicit Do-

2:7 Deliver thyself, O Zion, that dwellest with the daugh- minus 2:7 o Sion fuge quae habitas apud filiam Babylonis ter of Babylon.

2:8 For thus saith the LORD of hosts; After the glory hath 2:8 quia haec dicit Dominus exercituum post gloriam misit

he sent me unto the nations which spoiled you: for he that toucheth you toucheth the apple of his eye.

me ad gentes quae spoliauerunt uos qui enim tetigerit uos

2:9 For, behold, I will shake mine hand upon them, and tangit pupillam oculi eius 2:9 quia ecce ego leuo manum

they shall be a spoil to their servants: and ye shall know that the LORD of hosts hath sent me.

meam super eos et erunt praedae his qui seruiebant sibi et

Prophetia Zachariae

cognoscetis quia Dominus exercituum misit me 2:10 lauda 2:10 Sing and rejoice, O daughter of Zion: for, lo, I come,

and I will dwell in the midst of thee, saith the LORD.

et laetare filia Sion quia ecce ego uenio et habitabo in medio tui ait Dominus 2:11 et adplicabuntur gentes multae ad 2:11 And many nations shall be joined to the LORD in

that day, and shall be my people: and I will dwell in the midst of thee, and thou shalt know that the LORD of hosts

Dominum in die illa et erunt mihi in populum et habitabo in hath sent me unto thee. medio tui et scies quia Dominus exercituum misit me ad te 2:12 et possidebit Dominus Iudam partem suam in terra sanc- 2:12 And the LORD shall inherit Judah his portion in the

holy land, and shall choose Jerusalem again.

tificata et eliget adhuc Hierusalem 2:13 sileat omnis caro a 2:13 Be silent, O all flesh, before the LORD: for he is

raised up out of his holy habitation.

facie Domini quia consurrexit de habitaculo sancto suo

3:1 et ostendit mihi Iesum sacerdotem magnum stantem 3:1 And he shewed me Joshua the high priest standing before the angel of the LORD, and Satan standing at his right hand to resist him.

coram angelo Domini et Satan stabat a dextris eius ut aduersaretur ei 3:2 et dixit Dominus ad Satan increpet Do- 3:2 And the LORD said unto Satan, The LORD rebuke

thee, O Satan; even the LORD that hath chosen Jerusalem rebuke thee: is not this a brand plucked out of the fire?

minus in te Satan et increpet Dominus in te qui elegit Hierusalem numquid non iste torris est erutus de igne 3:3 et 3:3 Now Joshua was clothed with filthy garments, and

stood before the angel.

Iesus erat indutus uestibus sordidis et stabat ante faciem angeli 3:4 qui respondit et ait ad eos qui stabant coram se 3:4 And he answered and spake unto those that stood be-

fore him, saying, Take away the filthy garments from him. And unto him he said, Behold, I have caused thine iniquity to pass from thee, and I will clothe thee with change of raiment.

dicens auferte uestimenta sordida ab eo et dixit ad eum ecce abstuli a te iniquitatem tuam et indui te mutatoriis 3:5 et 3:5 And I said, Let them set a fair mitre upon his head. So

they set a fair mitre upon his head, and clothed him with garments. And the angel of the LORD stood by.

dixit ponite cidarim mundam super caput eius et posuerunt cidarim mundam super caput eius et induerunt eum uestibus et angelus Domini stabat 3:6 et contestabatur angelus 3:6 And the angel of the LORD protested unto Joshua,

saying,

Domini Iesum dicens 3:7 haec dicit Dominus exercituum 3:7 Thus saith the LORD of hosts; If thou wilt walk in my

ways, and if thou wilt keep my charge, then thou shalt also judge my house, and shalt also keep my courts, and I will

si in uiis meis ambulaueris et custodiam meam custodieris give thee places to walk among these that stand by. tu quoque iudicabis domum meam et custodies atria mea et dabo tibi ambulantes de his qui nunc hic adsistunt 3:8 audi 3:8 Hear now, O Joshua the high priest, thou, and thy fel-

lows that sit before thee: for they are men wondered at: for, behold, I will bring forth my servant the BRANCH.

Iesu sacerdos magne tu et amici tui qui habitant coram te quia uiri portendentes sunt ecce enim ego adducam seruum

meum orientem 3:9 quia ecce lapis quem dedi coram Iesu 3:9 For behold the stone that I have laid before Joshua; upon one stone shall be seven eyes: behold, I will engrave the graving thereof, saith the LORD of hosts, and I will

super lapidem unum septem oculi sunt ecce ego celabo scul- remove the iniquity of that land in one day. pturam eius ait Dominus exercituum et auferam iniquitatem terrae illius in die una 3:10 in die illa dicit Dominus ex- 3:10 In that day, saith the LORD of hosts, shall ye call

every man his neighbour under the vine and under the fig tree.

ercituum uocabit uir amicum suum subter uineam et subter ficum

4:1 et reuersus est angelus qui loquebatur in me et sus- 4:1 And the angel that talked with me came again, and

waked me, as a man that is wakened out of his sleep.

Biblia Sacra Vulgata

4:2 And said unto me, What seest thou? And I said, I have citauit me quasi uirum qui suscitatur de somno suo 4:2 et

looked, and behold a candlestick all of gold, with a bowl upon the top of it, and his seven lamps thereon, and seven pipes to the seven lamps, which are upon the top thereof: dixit ad me quid tu uides et dixi uidi et ecce candelabrum

aureum totum et lampas eius super caput ipsius et septem lucernae eius super illud septem et septem infusoria lucer-

4:3 And two olive trees by it, one upon the right side of nis quae erant super caput illius 4:3 et duae oliuae super

the bowl, and the other upon the left side thereof.

4:4 So I answered and spake to the angel that talked with illud una a dextris lampadis et una a sinistris eius 4:4 et re-

me, saying, What are these, my lord?

spondi et aio ad angelum qui loquebatur in me dicens quid

4:5 Then the angel that talked with me answered and said sunt haec domine mi 4:5 et respondit angelus qui loqueunto me, Knowest thou not what these be? And I said, No, my lord.

batur in me et dixit ad me numquid nescis quid sunt haec

4:6 Then he answered and spake unto me, saying, This is et dixi non domine mi 4:6 et respondit et ait ad me dicens

the word of the LORD unto Zerubbabel, saying, Not by might, nor by power, but by my spirit, saith the LORD of

hosts. hoc est uerbum Domini ad Zorobabel dicens non in exer-

citu nec in robore sed in spiritu meo dicit Dominus exerci-

4:7 Who art thou, O great mountain? before Zerubba- tuum 4:7 quis tu mons magne coram Zorobabel in planum

bel thou shalt become a plain: and he shall bring forth the headstone thereof with shoutings, crying, Grace, grace

unto it. et educet lapidem primarium et exaequabit gratiam gratiae 4:8 Moreover the word of the LORD came unto me, say- eius 4:8 et factum est uerbum Domini ad me dicens 4:9

ing, 4:9 The hands of Zerubbabel have laid the foundation of this house; his hands shall also finish it; and thou shalt know that the LORD of hosts hath sent me unto you.

manus Zorobabel fundauerunt domum istam et manus eius perficient eam et scietis quia Dominus exercituum misit me

4:10 For who hath despised the day of small things? for ad uos 4:10 quis enim despexit dies paruos et laetabuntur et

they shall rejoice, and shall see the plummet in the hand of Zerubbabel with those seven; they are the eyes of the

LORD, which run to and fro through the whole earth. uidebunt lapidem stagneum in manu Zorobabel septem isti 4:11 Then answered I, and said unto him, What are these oculi Domini qui discurrunt in uniuersa terra 4:11 et res-

two olive trees upon the right side of the candlestick and upon the left side thereof?

pondi et dixi ad eum quid sunt duae oliuae istae ad dextram

4:12 And I answered again, and said unto him, What be candelabri et ad sinistram eius 4:12 et respondi secundo et

these two olive branches which through the two golden pipes empty the golden oil out of themselves?

dixi ad eum quid sunt duae spicae oliuarum quae sunt iuxta

4:13 And he answered me and said, Knowest thou not duo rostra aurea in quibus sunt suffusoria ex auro 4:13 et ait

what these be? And I said, No, my lord.

ad me dicens numquid nescis quid sunt haec et dixi non do-

4:14 Then said he, These are the two anointed ones, that mine 4:14 et dixit isti duo filii olei qui adsistunt Dominatori

stand by the LORD of the whole earth.

uniuersae terrae

5:1 Then I turned, and lifted up mine eyes, and looked, 5:1 et conuersus sum et leuaui oculos meos et uidi et ecce

and behold a flying roll.

5:2 And he said unto me, What seest thou? And I an- uolumen uolans 5:2 et dixit ad me quid tu uides et dixi ego

swered, I see a flying roll; the length thereof is twenty cubits, and the breadth thereof ten cubits.

uideo uolumen uolans longitudo eius uiginti cubitorum et la-

5:3 Then said he unto me, This is the curse that goeth forth titudo eius decem cubitorum 5:3 et dixit ad me haec est maover the face of the whole earth: for every one that stealeth shall be cut off as on this side according to it; and every one that sweareth shall be cut off as on that side according to it.

ledictio quae egreditur super faciem omnis terrae quia omnis fur sicut ibi scriptum est iudicabitur et omnis iurans ex hoc

Prophetia Zachariae

similiter iudicabitur 5:4 educam illud dicit Dominus exer- 5:4 I will bring it forth, saith the LORD of hosts, and it

shall enter into the house of the thief, and into the house of him that sweareth falsely by my name: and it shall remain in the midst of his house, and shall consume it with the timber thereof and the stones thereof.

cituum et ueniet ad domum furis et ad domum iurantis in nomine meo mendaciter et commorabitur in medio domus eius et consumet eam et ligna eius et lapides eius 5:5 et egressus 5:5 Then the angel that talked with me went forth, and said

unto me, Lift up now thine eyes, and see what is this that goeth forth.

est angelus qui loquebatur in me et dixit ad me leua oculos tuos et uide quid est hoc quod egreditur 5:6 et dixi quidnam 5:6 And I said, What is it? And he said, This is an ephah

that goeth forth. He said moreover, This is their resemblance through all the earth.

est et ait haec est amphora egrediens et dixit haec est oculus eorum in uniuersa terra 5:7 et ecce talentum plumbi porta- 5:7 And, behold, there was lifted up a talent of lead: and

this is a woman that sitteth in the midst of the ephah.

batur et ecce mulier una sedens in medio amphorae 5:8 et 5:8 And he said, This is wickedness. And he cast it into

the midst of the ephah; and he cast the weight of lead upon the mouth thereof.

dixit haec est impietas et proiecit eam in medio amphorae et misit massam plumbeam in os eius 5:9 et leuaui oculos 5:9 Then lifted I up mine eyes, and looked, and, behold,

there came out two women, and the wind was in their wings; for they had wings like the wings of a stork: and

meos et uidi et ecce duae mulieres egredientes et spiritus in they lifted up the ephah between the earth and the heaven. alis earum et habebant alas quasi alas milui et leuauerunt

amphoram inter terram et caelum 5:10 et dixi ad angelum 5:10 Then said I to the angel that talked with me, Whither

do these bear the ephah?

qui loquebatur in me quo istae deferunt amphoram 5:11 et 5:11 And he said unto me, To build it an house in the land

of Shinar: and it shall be established, and set there upon her own base.

dixit ad me ut aedificetur ei domus in terra Sennaar et stabiliatur et ponatur ibi super basem suam

6:1 et conuersus sum et leuaui oculos meos et uidi et ecce 6:1 And I turned, and lifted up mine eyes, and looked, and, behold, there came four chariots out from between two mountains; and the mountains were mountains of brass.

quattuor quadrigae egredientes de medio duorum montium

et montes montes aerei 6:2 in quadriga prima equi rufi et 6:2 In the first chariot were red horses; and in the second

chariot black horses;

in quadriga secunda equi nigri 6:3 et in quadriga tertia equi 6:3 And in the third chariot white horses; and in the fourth

chariot grisled and bay horses.

albi et in quadriga quarta equi uarii fortes 6:4 et respondi et 6:4 Then I answered and said unto the angel that talked

with me, What are these, my lord?

dixi ad angelum qui loquebatur in me quid sunt haec domine mi 6:5 et respondit angelus et ait ad me isti sunt quattuor 6:5 And the angel answered and said unto me, These are

the four spirits of the heavens, which go forth from standing before the LORD of all the earth.

uenti caeli qui egrediuntur ut stent coram Dominatore omnis terrae 6:6 in quo erant equi nigri egrediebantur in terra 6:6 The black horses which are therein go forth into the

north country; and the white go forth after them; and the grisled go forth toward the south country.

aquilonis et albi egressi sunt post eos et uarii egressi sunt ad terram austri 6:7 qui autem erant robustissimi exierunt 6:7 And the bay went forth, and sought to go that they

might walk to and fro through the earth: and he said, Get you hence, walk to and fro through the earth. So they

et quaerebant ire et discurrere per omnem terram et dixit ite walked to and fro through the earth. perambulate terram et perambulauerunt terram 6:8 et uo- 6:8 Then cried he upon me, and spake unto me, saying,

Behold, these that go toward the north country have quieted my spirit in the north country.

cauit me et locutus est ad me dicens ecce qui egrediuntur in terram aquilonis requiescere fecerunt spiritum meum in terra aquilonis 6:9 et factum est uerbum Domini ad me di- 6:9 And the word of the LORD came unto me, saying,

Biblia Sacra Vulgata

6:10 Take of them of the captivity, even of Heldai, of To- cens 6:10 sume a transmigratione ab Oldai et a Tobia et ab bijah, and of Jedaiah, which are come from Babylon, and come thou the same day, and go into the house of Josiah

the son of Zephaniah; Idaia et uenies tu in die illa et intrabis domum Iosiae filii So-

6:11 Then take silver and gold, and make crowns, and set foniae qui uenerunt de Babylone 6:11 et sumes argentum them upon the head of Joshua the son of Josedech, the high priest;

et aurum et facies coronas et pones in capite Iesu filii Iose-

6:12 And speak unto him, saying, Thus speaketh the dech sacerdotis magni 6:12 et loqueris ad eum dicens haec

LORD of hosts, saying, Behold the man whose name is The BRANCH; and he shall grow up out of his place, and

he shall build the temple of the LORD: ait Dominus exercituum dicens ecce uir Oriens nomen eius 6:13 Even he shall build the temple of the LORD; and he et subter eum orietur et aedificabit templum Domino 6:13

shall bear the glory, and shall sit and rule upon his throne; and he shall be a priest upon his throne: and the counsel

of peace shall be between them both. et ipse extruet templum Domino et ipse portabit gloriam et

sedebit et dominabitur super solio suo et erit sacerdos su-

6:14 And the crowns shall be to Helem, and to Tobijah, per solio suo et consilium pacis erit inter duos illos 6:14 et

and to Jedaiah, and to Hen the son of Zephaniah, for a memorial in the temple of the LORD.

coronae erunt Helem et Tobiae et Idaiae et Hen filio Sofo-

6:15 And they that are far off shall come and build in the niae memoriale in templo Domini 6:15 et qui procul sunt

temple of the LORD, and ye shall know that the LORD of hosts hath sent me unto you. And this shall come to pass,

if ye will diligently obey the voice of the LORD your God. uenient et aedificabunt in templo Domini et scietis quia Dominus exercituum misit me ad uos erit autem hoc si auditu audieritis uocem Domini Dei uestri

7:1 And it came to pass in the fourth year of king Darius, 7:1 et factum est in anno quarto Darii regis factum est that the word of the LORD came unto Zechariah in the fourth day of the ninth month, even in Chisleu;

uerbum Domini ad Zacchariam in quarta mensis noni qui

7:2 When they had sent unto the house of God Sherezer est casleu 7:2 et miserunt ad domum Dei Sarasar et Ro-

and Regemmelech, and their men, to pray before the LORD,

gomelech et uiri qui erant cum eo ad deprecandam faciem

7:3 And to speak unto the priests which were in the house Domini 7:3 ut dicerent sacerdotibus domus Domini exer-

of the LORD of hosts, and to the prophets, saying, Should I weep in the fifth month, separating myself, as I have done

these so many years? cituum et prophetis loquentes numquid flendum mihi est in

mense quinto uel sanctificare me debeo sicuti feci iam mul-

7:4 Then came the word of the LORD of hosts unto me, tis annis 7:4 et factum est uerbum Domini exercituum ad

saying, 7:5 Speak unto all the people of the land, and to the priests, me dicens 7:5 loquere ad omnem populum terrae et ad sasaying, When ye fasted and mourned in the fifth and seventh month, even those seventy years, did ye at all fast

unto me, even to me? cerdotes dicens cum ieiunaretis et plangeretis in quinto et

septimo per hos septuaginta annos numquid ieiunium ieiu-

7:6 And when ye did eat, and when ye did drink, did not nastis mihi 7:6 et cum comedistis et cum bibistis numquid

ye eat for yourselves, and drink for yourselves?

7:7 Should ye not hear the words which the LORD hath non uobis comedistis et uobismet ipsis bibistis 7:7 num-

cried by the former prophets, when Jerusalem was inhabited and in prosperity, and the cities thereof round about

her, when men inhabited the south and the plain? quid non sunt uerba quae locutus est Dominus in manu prophetarum priorum cum adhuc Hierusalem habitaretur et esset opulenta ipsa et urbes in circuitu eius et ad austrum et in

7:8 And the word of the LORD came unto Zechariah, say- campestribus habitaretur 7:8 et factum est uerbum Domini

ing,

7:9 Thus speaketh the LORD of hosts, saying, Execute ad Zacchariam dicens 7:9 haec ait Dominus exercituum di-

true judgment, and shew mercy and compassions every man to his brother:

Prophetia Zachariae

cens iudicium uerum iudicate et misericordiam et miserationes facite unusquisque cum fratre suo 7:10 et uiduam et 7:10 And oppress not the widow, nor the fatherless, the

stranger, nor the poor; and let none of you imagine evil against his brother in your heart.

pupillum et aduenam et pauperem nolite calumniari et malum uir fratri suo non cogitet in corde suo 7:11 et noluerunt 7:11 But they refused to hearken, and pulled away the

shoulder, and stopped their ears, that they should not hear.

adtendere et uerterunt scapulam recedentem et aures suas adgrauauerunt ne audirent 7:12 et cor suum posuerunt ad- 7:12 Yea, they made their hearts as an adamant stone, lest

they should hear the law, and the words which the LORD of hosts hath sent in his spirit by the former prophets:

amantem ne audirent legem et uerba quae misit Dominus therefore came a great wrath from the LORD of hosts. exercituum in spiritu suo per manum prophetarum priorum

et facta est indignatio magna a Domino exercituum 7:13 et 7:13 Therefore it is come to pass, that as he cried, and they would not hear; so they cried, and I would not hear, saith the LORD of hosts:

factum est sicut locutus est et non audierunt sic clamabunt et non exaudiam dicit Dominus exercituum 7:14 et dispersi 7:14 But I scattered them with a whirlwind among all the

nations whom they knew not. Thus the land was desolate after them, that no man passed through nor returned: for

eos per omnia regna quae nesciunt et terra desolata est ab eis they laid the pleasant land desolate. eo quod non esset transiens et reuertens et posuerunt terram desiderabilem in desertum

8:1 et factum est uerbum Domini exercituum dicens 8:2 8:1 Again the word of the LORD of hosts came to me,

saying,

8:2 Thus saith the LORD of hosts; I was jealous for Zion with great jealousy, and I was jealous for her with great fury.

haec dicit Dominus exercituum zelatus sum Sion zelo magno et indignatione magna zelatus sum eam 8:3 haec dicit Do- _{8:3 Thus saith the LORD; I am returned unto Zion, and}

will dwell in the midst of Jerusalem: and Jerusalem shall be called a city of truth; and the mountain of the LORD of

minus exercituum reuersus sum ad Sion et habitabo in me- _{hosts the holy mountain.} dio Hierusalem et uocabitur Hierusalem ciuitas ueritatis et mons Domini exercituum mons sanctificatus 8:4 haec di- 8:4 Thus saith the LORD of hosts; There shall yet old men

and old women dwell in the streets of Jerusalem, and every man with his staff in his hand for very age.

cit Dominus exercituum adhuc habitabunt senes et anus in plateis Hierusalem et uiri baculus in manu eius prae multitudine dierum 8:5 et plateae ciuitatis conplebuntur infanti- 8:5 And the streets of the city shall be full of boys and girls

playing in the streets thereof.

bus et puellis ludentibus in plateis eius 8:6 haec dicit Do- 8:6 Thus saith the LORD of hosts; If it be marvellous in

the eyes of the remnant of this people in these days, should it also be marvellous in mine eyes? saith the LORD of

minus exercituum si difficile uidebitur in oculis reliquiarum hosts. populi huius in diebus illis numquid in oculis meis difficile erit dicit Dominus exercituum 8:7 haec dicit Dominus ex- 8:7 Thus saith the LORD of hosts; Behold, I will save my

people from the east country, and from the west country;

ercituum ecce ego saluabo populum meum de terra orientis et de terra occasus solis 8:8 et adducam eos et habitabunt 8:8 And I will bring them, and they shall dwell in the midst

of Jerusalem: and they shall be my people, and I will be their God, in truth and in righteousness.

in medio Hierusalem et erunt mihi in populum et ego ero eis in Deum in ueritate et iustitia 8:9 haec dicit Dominus 8:9 Thus saith the LORD of hosts; Let your hands be

strong, ye that hear in these days these words by the mouth of the prophets, which were in the day that the foundation of the house of the LORD of hosts was laid, that the temple might be built.

exercituum confortentur manus uestrae qui auditis in diebus his sermones istos per os prophetarum in die qua fun-

Biblia Sacra Vulgata

data est domus Domini exercituum ut templum aedificare-

8:10 For before these days there was no hire for man, nor tur 8:10 siquidem ante dies illos merces hominum non erat

any hire for beast; neither was there any peace to him that went out or came in because of the affliction: for I set all

men every one against his neighbour. nec merces iumentorum erat neque introeunti et exeunti erat

pax prae tribulatione et dimisi omnes homines unumquem-

8:11 But now I will not be unto the residue of this people que contra proximum suum 8:11 nunc autem non iuxta dies as in the former days, saith the LORD of hosts.

priores ego faciam reliquiis populi huius dicit Dominus exer-

8:12 For the seed shall be prosperous; the vine shall give cituum 8:12 sed semen pacis erit uinea dabit fructum suum her fruit, and the ground shall give her increase, and the heavens shall give their dew; and I will cause the remnant

of this people to possess all these things. et terra dabit germen suum et caeli dabunt rorem suum et 8:13 And it shall come to pass, that as ye were a curse possidere faciam reliquias populi huius uniuersa haec 8:13

among the heathen, O house of Judah, and house of Israel; so will I save you, and ye shall be a blessing: fear not, but

let your hands be strong. et erit sicut eratis maledictio in gentibus domus Iuda et do-

mus Israhel sic saluabo uos et eritis benedictio nolite timere

8:14 For thus saith the LORD of hosts; As I thought to confortentur manus uestrae 8:14 quia haec dicit Dominus

punish you, when your fathers provoked me to wrath, saith the LORD of hosts, and I repented not:

exercituum sicut cogitaui ut adfligerem uos cum ad iracun-

8:15 So again have I thought in these days to do well unto diam prouocassent patres uestri me dicit Dominus 8:15 et

Jerusalem and to the house of Judah: fear ye not.

non sum misertus sic conuersus cogitaui in diebus istis ut be-

8:16 These are the things that ye shall do; Speak ye every nefaciam Hierusalem et domui Iuda nolite timere 8:16 haec

man the truth to his neighbour; execute the judgment of truth and peace in your gates:

sunt ergo uerba quae facietis loquimini ueritatem unusquisque cum proximo suo ueritatem et iudicium pacis iudicate

8:17 And let none of you imagine evil in your hearts in portis uestris 8:17 et unusquisque malum contra amicum

against his neighbour; and love no false oath: for all these are things that I hate, saith the LORD.

suum ne cogitetis in cordibus uestris et iuramentum mendax ne diligatis omnia enim haec sunt quae odi dicit Domi-

8:18 And the word of the LORD of hosts came unto me, nus 8:18 et factum est uerbum Domini exercituum ad me

saying,

8:19 Thus saith the LORD of hosts; The fast of the fourth dicens 8:19 haec dicit Dominus exercituum ieiunium quarti

month, and the fast of the fifth, and the fast of the seventh, and the fast of the tenth, shall be to the house of Judah joy and gladness, and cheerful feasts; therefore love the truth and peace.

et ieiunium quinti et ieiunium septimi et ieiunium decimi erit domui Iuda in gaudium et in laetitiam et in sollemnita-

8:20 Thus saith the LORD of hosts; It shall yet come to tes praeclaras ueritatem tantum et pacem diligite 8:20 haec

pass, that there shall come people, and the inhabitants of many cities:

dicit Dominus exercituum usquequo ueniant populi et ha-

8:21 And the inhabitants of one city shall go to another, bitent in ciuitatibus multis 8:21 et uadant habitatores unus

saying, Let us go speedily to pray before the LORD, and to seek the LORD of hosts: I will go also.

ad alterum dicentes eamus et deprecemur faciem Domini et

8:22 Yea, many people and strong nations shall come to quaeramus Dominum exercituum uadam etiam ego 8:22 et seek the LORD of hosts in Jerusalem, and to pray before the LORD.

uenient populi multi et gentes robustae ad quaerendum Dominum exercituum in Hierusalem et deprecandam faciem

8:23 Thus saith the LORD of hosts; In those days it shall Domini 8:23 haec dicit Dominus exercituum in diebus il-

come to pass, that ten men shall take hold out of all languages of the nations, even shall take hold of the skirt of him that is a Jew, saying, We will go with you: for we have heard that God is with you.

Prophetia Zachariae

lis in quibus adprehendent decem homines ex omnibus linguis gentium et adprehendent fimbriam uiri iudaei dicentes ibimus uobiscum audiuimus enim quoniam Deus uobiscum est

9:1 onus uerbi Domini in terra Adrach et Damasci requiei 9:1 The burden of the word of the LORD in the land of

Hadrach, and Damascus shall be the rest thereof: when the eyes of man, as of all the tribes of Israel, shall be toward

eius quia Domini est oculus hominis et omnium tribuum Is- the LORD. rahel 9:2 Emath quoque in terminis eius et Tyrus et Sidon 9:2 And Hamath also shall border thereby; Tyrus, and Zi-

don, though it be very wise.

adsumpserunt quippe sibi sapientiam ualde 9:3 et aedifica- 9:3 And Tyrus did build herself a strong hold, and heaped

up silver as the dust, and fine gold as the mire of the streets.

uit Tyrus munitionem suam et coaceruauit argentum quasi

humum et aurum ut lutum platearum 9:4 ecce Dominus 9:4 Behold, the LORD will cast her out, and he will smite

her power in the sea; and she shall be devoured with fire.

possidebit eam et percutiet in mari fortitudinem eius et haec igni deuorabitur 9:5 uidebit Ascalon et timebit et Gaza et 9:5 Ashkelon shall see it, and fear; Gaza also shall see

it, and be very sorrowful, and Ekron; for her expectation shall be ashamed; and the king shall perish from Gaza, and

dolebit nimis et Accaron quoniam confusa est spes eius et Ashkelon shall not be inhabited. peribit rex de Gaza et Ascalon non habitabitur 9:6 et sede- 9:6 And a bastard shall dwell in Ashdod, and I will cut off

the pride of the Philistines.

bit separator in Azoto et disperdam superbiam Philisthinorum 9:7 et auferam sanguinem eius de ore eius et abomina- 9:7 And I will take away his blood out of his mouth, and

his abominations from between his teeth: but he that remaineth, even he, shall be for our God, and he shall be as

tiones eius de medio dentium eius et relinquetur etiam ipse a governor in Judah, and Ekron as a Jebusite. Deo nostro et erit quasi dux in Iuda et Accaron quasi Iebuseus 9:8 et circumdabo domum meam ex his qui militant 9:8 And I will encamp about mine house because of the

army, because of him that passeth by, and because of him that returneth: and no oppressor shall pass through them

mihi euntes et reuertentes et non transibit super eos ultra ex- any more: for now have I seen with mine eyes. actor quia nunc uidi in oculis meis 9:9 exulta satis filia Sion 9:9 Rejoice greatly, O daughter of Zion; shout, O daughter

of Jerusalem: behold, thy King cometh unto thee: he is just, and having salvation; lowly, and riding upon an ass,

iubila filia Hierusalem ecce rex tuus ueniet tibi iustus et sal- and upon a colt the foal of an ass. uator ipse pauper et ascendens super asinum et super pullum filium asinae 9:10 et disperdam quadrigam ex Ephraim et 9:10 And I will cut off the chariot from Ephraim, and the

horse from Jerusalem, and the battle bow shall be cut off: and he shall speak peace unto the heathen: and his dominion shall be from sea even to sea, and from the river even to the ends of the earth.

equum de Hierusalem et dissipabitur arcus belli et loquetur pacem gentibus et potestas eius a mari usque ad mare et

a fluminibus usque ad fines terrae 9:11 tu quoque in san- 9:11 As for thee also, by the blood of thy covenant I have

sent forth thy prisoners out of the pit wherein is no water.

guine testamenti tui emisisti uinctos tuos de lacu in quo non est aqua 9:12 conuertimini ad munitionem uincti spei hodie 9:12 Turn you to the strong hold, ye prisoners of hope:

even to day do I declare that I will render double unto thee;

quoque adnuntians duplicia reddam tibi 9:13 quoniam ex-

9:13 When I have bent Judah for me, filled the bow with Ephraim, and raised up thy sons, O Zion, against thy sons, O Greece, and made thee as the sword of a mighty man.

tendi mihi Iudam quasi arcum impleui Ephraim et suscitabo filios tuos Sion super filios tuos Graecia et ponam te quasi gladium fortium 9:14 et Dominus Deus super eos uidebi- 9:14 And the LORD shall be seen over them, and his ar-

row shall go forth as the lightning: and the LORD God shall blow the trumpet, and shall go with whirlwinds of the south.

Biblia Sacra Vulgata

tur et exibit ut fulgur iaculum eius et Dominus Deus in tuba

9:15 The LORD of hosts shall defend them; and they shall canet et uadet in turbine austri 9:15 Dominus exercituum

devour, and subdue with sling stones; and they shall drink, and make a noise as through wine; and they shall be filled

like bowls, and as the corners of the altar. proteget eos et deuorabunt et subicient lapidibus fundae et

bibentes inebriabuntur quasi uino et replebuntur ut fialae et

9:16 And the LORD their God shall save them in that day quasi cornua altaris 9:16 et saluabit eos Dominus Deus eo-

as the flock of his people: for they shall be as the stones of a crown, lifted up as an ensign upon his land.

rum in die illa ut gregem populi sui quia lapides sancti ele-

9:17 For how great is his goodness, and how great is his uantur super terram eius 9:17 quid enim bonum eius est et

beauty! corn shall make the young men cheerful, and new wine the maids.

quid pulchrum eius nisi frumentum electorum et uinum germinans uirgines

10:1 Ask ye of the LORD rain in the time of the latter 10:1 petite a Domino pluuiam in tempore serotino et Do-

rain; so the LORD shall make bright clouds, and give them showers of rain, to every one grass in the field.

minus faciet niues et pluuiam imbris dabit eis singulis her-

10:2 For the idols have spoken vanity, and the diviners bam in agro 10:2 quia simulacra locuta sunt inutile et di-

have seen a lie, and have told false dreams; they comfort in vain: therefore they went their way as a flock, they were

troubled, because there was no shepherd. uini uiderunt mendacium et somniatores frustra locuti sunt

uane consolabantur idcirco abducti sunt quasi grex adfligen-

10:3 Mine anger was kindled against the shepherds, and tur quia non est eis pastor 10:3 super pastores iratus est

I punished the goats: for the LORD of hosts hath visited his flock the house of Judah, and hath made them as his

goodly horse in the battle. furor meus et super hircos uisitabo quia uisitauit Dominus

exercituum gregem suum domum Iuda et posuit eos quasi

10:4 Out of him came forth the corner, out of him the nail, equum gloriae suae in bello 10:4 ex ipso angulus ex ipso out of him the battle bow, out of him every oppressor together.

paxillus ex ipso arcus proelii ex ipso egredietur omnis exac-

10:5 And they shall be as mighty men, which tread down tor simul 10:5 et erunt quasi fortes conculcantes lutum uia-

their enemies in the mire of the streets in the battle: and they shall fight, because the LORD is with them, and the

riders on horses shall be confounded. rum in proelio et bellabunt quia Dominus cum eis et confun-10:6 And I will strengthen the house of Judah, and I will dentur ascensores equorum 10:6 et confortabo domum Iuda

save the house of Joseph, and I will bring them again to place them; for I have mercy upon them: and they shall be as though I had not cast them off: for I am the LORD their God, and will hear them.

et domum Ioseph saluabo et conuertam eos quia miserebor eorum et erunt sicut fuerunt quando non proieceram eos ego

10:7 And they of Ephraim shall be like a mighty man, and enim Dominus Deus eorum et exaudiam eos 10:7 et erunt their heart shall rejoice as through wine: yea, their children shall see it, and be glad; their heart shall rejoice in

the LORD. quasi fortes Ephraim et laetabitur cor eorum quasi a uino et

filii eorum uidebunt et laetabuntur et exultabit cor eorum in

10:8 I will hiss for them, and gather them; for I have Domino 10:8 sibilabo eis et congregabo illos quia redeei

redeemed them: and they shall increase as they have increased.

eos et multiplicabo eos sicut ante fuerant multiplicati 10:9

10:9 And I will sow them among the people: and they shall remember me in far countries; and they shall live with their children, and turn again.

et seminabo eos in populis et de longe recordabuntur mei et

10:10 I will bring them again also out of the land of Egypt, uiuent cum filiis suis et reuertentur 10:10 et reducam eos and gather them out of Assyria; and I will bring them into the land of Gilead and Lebanon; and place shall not be

found for them. de terra Aegypti et de Assyriis congregabo eos et ad terram Galaad et Libani adducam eos et non inuenietur eis locus

Prophetia Zachariae

10:11 et transibit in maris freto et percutiet in mari fluctus et 10:11 And he shall pass through the sea with affliction,

and shall smite the waves in the sea, and all the deeps of the river shall dry up: and the pride of Assyria shall be

confundentur omnia profunda Fluminis et humiliabitur su- brought down, and the sceptre of Egypt shall depart away. perbia Assur et sceptrum Aegypti recedet 10:12 confortabo 10:12 And I will strengthen them in the LORD; and they

shall walk up and down in his name, saith the LORD.

eos in Domino et in nomine eius ambulabunt dicit Dominus

11:1 aperi Libane portas tuas et comedat ignis cedros tuas 11:1 Open thy doors, O Lebanon, that the fire may devour

thy cedars.

11:2 ulula abies quia cecidit cedrus quoniam magnifici ua- 11:2 Howl, fir tree; for the cedar is fallen; because the

mighty are spoiled: howl, O ye oaks of Bashan; for the forest of the vintage is come down.

stati sunt ululate quercus Basan quoniam succisus est saltus munitus 11:3 uox ululatus pastorum quia uastata est 11:3 There is a voice of the howling of the shepherds; for

their glory is spoiled: a voice of the roaring of young lions; for the pride of Jordan is spoiled.

magnificentia eorum uox rugitus leonum quoniam uastata est superbia Iordanis 11:4 haec dicit Dominus Deus meus 11:4 Thus saith the LORD my God; Feed the flock of the

slaughter;

pasce pecora occisionis 11:5 quae qui possederant occide- 11:5 Whose possessors slay them, and hold themselves not

guilty: and they that sell them say, Blessed be the LORD; for I am rich: and their own shepherds pity them not.

bant et non dolebant et uendebant ea dicentes benedictus Dominus diuites facti sumus et pastores eorum non parcebant eis 11:6 et ego non parcam ultra super habitantes terram 11:6 For I will no more pity the inhabitants of the land,

saith the LORD: but, lo, I will deliver the men every one into his neighbour’s hand, and into the hand of his king: and they shall smite the land, and out of their hand I will not deliver them.

dicit Dominus ecce ego tradam homines unumquemque in manu proximi sui et in manu regis sui et concident terram et non eruam de manu eorum 11:7 et pascam pecus occisionis 11:7 And I will feed the flock of slaughter, even you, O

poor of the flock. And I took unto me two staves; the one I called Beauty, and the other I called Bands; and I fed the

propter hoc o pauperes gregis et adsumpsi mihi duas uirgas flock. unam uocaui Decorem et alteram uocaui Funiculos et paui gregem 11:8 et succidi tres pastores in mense uno et con- 11:8 Three shepherds also I cut off in one month; and my

soul lothed them, and their soul also abhorred me.

tracta est anima mea in eis siquidem anima eorum uariauit

in me 11:9 et dixi non pascam uos quod moritur moriatur 11:9 Then said I, I will not feed you: that that dieth, let it die; and that that is to be cut off, let it be cut off; and let the rest eat every one the flesh of another.

et quod succiditur succidatur et reliqui uorent unusquisque

carnem proximi sui 11:10 et tuli uirgam meam quae uoca- 11:10 And I took my staff, even Beauty, and cut it asunder, that I might break my covenant which I had made with all the people.

batur Decus et abscidi eam ut irritum facerem foedus meum quod percussi cum omnibus populis 11:11 et in irritum de- 11:11 And it was broken in that day: and so the poor of

the flock that waited upon me knew that it was the word of the LORD.

ductum est in die illa et cognouerunt sic pauperes gregis qui custodiunt mihi quia uerbum Domini est 11:12 et dixi ad 11:12 And I said unto them, If ye think good, give me my

price; and if not, forbear. So they weighed for my price thirty pieces of silver.

eos si bonum est in oculis uestris adferte mercedem meam et si non quiescite et adpenderunt mercedem meam triginta argenteos 11:13 et dixit Dominus ad me proice illud ad sta- 11:13 And the LORD said unto me, Cast it unto the potter:

a goodly price that I was prised at of them. And I took the thirty pieces of silver, and cast them to the potter in the

tuarium decorum pretium quod adpretiatus sum ab eis et tuli house of the LORD. triginta argenteos et proieci illos in domo Domini ad statua-

Biblia Sacra Vulgata

11:14 Then I cut asunder mine other staff, even Bands, that rium 11:14 et praecidi uirgam meam secundam quae appel-I might break the brotherhood between Judah and Israel.

labatur Funiculus ut dissoluerem germanitatem inter Iudam

11:15 And the LORD said unto me, Take unto thee yet the et inter Israhel 11:15 et dixit Dominus ad me adhuc sume

instruments of a foolish shepherd.

11:16 For, lo, I will raise up a shepherd in the land, which tibi uasa pastoris stulti 11:16 quia ecce ego suscitabo pasto-

shall not visit those that be cut off, neither shall seek the young one, nor heal that that is broken, nor feed that that standeth still: but he shall eat the flesh of the fat, and tear their claws in pieces.

rem in terra qui derelicta non uisitabit dispersum non quaeret et contritum non sanabit et id quod stat non enutriet et carnes

11:17 Woe to the idol shepherd that leaveth the flock! the pinguium comedet et ungulas eorum dissoluet 11:17 o pasword shall be upon his arm, and upon his right eye: his arm shall be clean dried up, and his right eye shall be ut-

terly darkened. stor et idolum derelinquens gregem gladius super brachium eius et super oculum dextrum eius brachium eius ariditate siccabitur et oculus dexter eius tenebrescens obscurabitur

12:1 The burden of the word of the LORD for Israel, saith 12:1 onus uerbi Domini super Israhel dixit Dominus exten-

the LORD, which stretcheth forth the heavens, and layeth the foundation of the earth, and formeth the spirit of man

within him. dens caelum et fundans terram et fingens spiritum hominis in 12:2 Behold, I will make Jerusalem a cup of trembling eo 12:2 ecce ego ponam Hierusalem superliminare crapu-

unto all the people round about, when they shall be in the siege both against Judah and against Jerusalem.

lae omnibus populis in circuitu sed et Iuda erit in obsidione

12:3 And in that day will I make Jerusalem a burdensome contra Hierusalem 12:3 et erit in die illa ponam Hierusalem

stone for all people: all that burden themselves with it shall be cut in pieces, though all the people of the earth be gath-

ered together against it. lapidem oneris cunctis populis omnes qui leuabunt eam con-

cisione lacerabuntur et colligentur aduersum eam omnia re-

12:4 In that day, saith the LORD, I will smite every horse gna terrae 12:4 in die illa dicit Dominus percutiam omnem

with astonishment, and his rider with madness: and I will open mine eyes upon the house of Judah, and will smite

every horse of the people with blindness. equum in stuporem et ascensorem eius in amentiam et super

domum Iuda aperiam oculos meos et omnem equum popu-

12:5 And the governors of Judah shall say in their heart, lorum percutiam in caecitate 12:5 et dicent duces Iuda in

The inhabitants of Jerusalem shall be my strength in the LORD of hosts their God.

corde suo confortentur mihi habitatores Hierusalem in Do-

12:6 In that day will I make the governors of Judah like an mino exercituum Deo eorum 12:6 in die illo ponam duces hearth of fire among the wood, and like a torch of fire in a sheaf; and they shall devour all the people round about, on the right hand and on the left: and Jerusalem shall be inhabited again in her own place, even in Jerusalem.

Iuda sicut caminum ignis in lignis et sicut facem ignis in faeno et deuorabunt ad dextram et ad sinistram omnes populos in circuitu et habitabitur Hierusalem rursum in loco suo

12:7 The LORD also shall save the tents of Judah first, that in Hierusalem 12:7 et saluabit Dominus tabernacula Iuda the glory of the house of David and the glory of the inhabitants of Jerusalem do not magnify themselves against

Judah. sicut in principio ut non magnifice glorietur domus Dauid 12:8 In that day shall the LORD defend the inhabitants of et gloria habitantium Hierusalem contra Iudam 12:8 in die

Jerusalem; and he that is feeble among them at that day shall be as David; and the house of David shall be as God,

as the angel of the LORD before them. illo proteget Dominus habitatores Hierusalem et erit qui of-

fenderit ex eis in die illa quasi Dauid et domus Dauid quasi

12:9 And it shall come to pass in that day, that I will seek Dei sicut angelus Domini in conspectu eius 12:9 et erit in

to destroy all the nations that come against Jerusalem.

die illa quaeram conterere omnes gentes quae ueniunt con-

Prophetia Zachariae

tra Hierusalem 12:10 et effundam super domum Dauid et 12:10 And I will pour upon the house of David, and upon

the inhabitants of Jerusalem, the spirit of grace and of supplications: and they shall look upon me whom they have pierced, and they shall mourn for him, as one mourneth for his only son, and shall be in bitterness for him, as one that is in bitterness for his firstborn.

super habitatores Hierusalem spiritum gratiae et precum et aspicient ad me quem confixerunt et plangent eum planctu quasi super unigenitum et dolebunt super eum ut doleri solet in morte primogeniti 12:11 in die illa magnus erit planctus 12:11 In that day shall there be a great mourning in

Jerusalem, as the mourning of Hadadrimmon in the valley of Megiddon.

in Hierusalem sicut planctus Adadremmon in campo Mageddon 12:12 et planget terra familiae et familiae seorsum 12:12 And the land shall mourn, every family apart; the

family of the house of David apart, and their wives apart; the family of the house of Nathan apart, and their wives

familiae domus Dauid seorsum et mulieres eorum seorsum apart; 12:13 familiae domus Nathan seorsum et mulieres eorum se- 12:13 The family of the house of Levi apart, and their

wives apart; the family of Shimei apart, and their wives apart;

orsum familiae domus Leui seorsum et mulieres eorum seorsum familiae Semei seorsum et mulieres eorum seorsum 12:14 omnes familiae reliquae familiae et familiae seorsum 12:14 All the families that remain, every family apart, and

their wives apart.

et mulieres eorum seorsum

13:1 in die illa erit fons patens domus Dauid et habitan- 13:1 In that day there shall be a fountain opened to the

house of David and to the inhabitants of Jerusalem for sin and for uncleanness.

tibus Hierusalem in ablutionem peccatoris et menstruatae 13:2 et erit in die illa dicit Dominus exercituum disperdam 13:2 And it shall come to pass in that day, saith the LORD

of hosts, that I will cut off the names of the idols out of the land, and they shall no more be remembered: and also I will cause the prophets and the unclean spirit to pass out of the land.

nomina idolorum de terra et non memorabuntur ultra et prophetas et spiritum inmundum auferam de terra 13:3 et erit 13:3 And it shall come to pass, that when any shall yet

prophesy, then his father and his mother that begat him shall say unto him, Thou shalt not live; for thou speakest lies in the name of the LORD: and his father and his mother that begat him shall thrust him through when he prophesieth.

cum prophetauerit quispiam ultra dicent ei pater eius et mater eius qui genuerunt eum non uiues quia mendacium locutus es in nomine Domini et configent eum pater eius et mater eius genitores eius cum prophetauerit 13:4 et erit in 13:4 And it shall come to pass in that day, that the prophets

shall be ashamed every one of his vision, when he hath prophesied; neither shall they wear a rough garment to de-

die illa confundentur prophetae unusquisque ex uisione sua ceive: cum prophetauerit nec operientur pallio saccino ut mentiantur 13:5 sed dicet non sum propheta homo agricola ego sum 13:5 But he shall say, I am no prophet, I am an husband-

man; for man taught me to keep cattle from my youth.

quoniam Adam exemplum meum ab adulescentia mea 13:6 13:6 And one shall say unto him, What are these wounds

in thine hands? Then he shall answer, Those with which I was wounded in the house of my friends.

et dicetur ei quid sunt plagae istae in medio manuum tuarum et dicet his plagatus sum in domo eorum qui diligebant me 13:7 framea suscitare super pastorem meum et super ui- 13:7 Awake, O sword, against my shepherd, and against

the man that is my fellow, saith the LORD of hosts: smite the shepherd, and the sheep shall be scattered: and I will

rum coherentem mihi dicit Dominus exercituum percute pa- turn mine hand upon the little ones. storem et dispergantur oues et conuertam manum meam ad paruulos 13:8 et erunt in omni terra dicit Dominus partes 13:8 And it shall come to pass, that in all the land, saith

the LORD, two parts therein shall be cut off and die; but the third shall be left therein.

duae in ea disperdentur et deficient et tertia pars relinque-

Biblia Sacra Vulgata

13:9 And I will bring the third part through the fire, and tur in ea 13:9 et ducam tertiam partem per ignem et uram will refine them as silver is refined, and will try them as gold is tried: they shall call on my name, and I will hear them: I will say, It is my people: and they shall say, The LORD is my God.

eas sicut uritur argentum et probabo eos sicut probatur aurum ipse uocabit nomen meum et ego exaudiam eum dicam populus meus es et ipse dicet Dominus Deus meus

14:1 Behold, the day of the LORD cometh, and thy spoil 14:1 ecce dies ueniunt Domini et diuidentur spolia tua in

shall be divided in the midst of thee. 14:2 For I will gather all nations against Jerusalem to bat- medio tui 14:2 et congregabo omnes gentes ad Hierusalem tle; and the city shall be taken, and the houses rifled, and the women ravished; and half of the city shall go forth into captivity, and the residue of the people shall not be cut off from the city.

in proelium et capietur ciuitas et uastabuntur domus et mulieres uiolabuntur et egredietur media pars ciuitatis in cap-

14:3 Then shall the LORD go forth, and fight against those tiuitatem et reliquum populi non auferetur ex urbe 14:3 et

nations, as when he fought in the day of battle.

egredietur Dominus et proeliabitur contra gentes illas sicut

14:4 And his feet shall stand in that day upon the mount proeliatus est in die certaminis 14:4 et stabunt pedes eius

of Olives, which is before Jerusalem on the east, and the mount of Olives shall cleave in the midst thereof toward the east and toward the west, and there shall be a very great valley; and half of the mountain shall remove toward the north, and half of it toward the south.

in die illa super montem Oliuarum qui est contra Hierusalem ad orientem et scindetur mons Oliuarum ex media parte sui ad orientem et occidentem praerupto grandi ualde et separabitur medium montis ad aquilonem et medium eius ad

14:5 And ye shall flee to the valley of the mountains; for meridiem 14:5 et fugietis ad uallem montium meorum quo-

the valley of the mountains shall reach unto Azal: yea, ye shall flee, like as ye fled from before the earthquake in the days of Uzziah king of Judah: and the LORD my God shall come, and all the saints with thee.

niam coniungetur uallis montium usque ad proximum et fugietis sicut fugistis a facie terraemotus in diebus Oziae regis Iuda et ueniet Dominus Deus meus omnesque sancti cum eo

14:6 And it shall come to pass in that day, that the light 14:6 et erit in die illa non erit lux sed frigus et gelu 14:7 et

shall not be clear, nor dark: 14:7 But it shall be one day which shall be known to the LORD, not day, nor night: but it shall come to pass, that at evening time it shall be light.

erit dies una quae nota est Domino non dies neque nox et

14:8 And it shall be in that day, that living waters shall go in tempore uesperae erit lux 14:8 et erit in die illa exibunt

out from Jerusalem; half of them toward the former sea, and half of them toward the hinder sea: in summer and in

winter shall it be. aquae uiuae de Hierusalem medium earum ad mare orientale

et medium earum ad mare nouissimum in aestate et in hieme

14:9 And the LORD shall be king over all the earth: in erunt 14:9 et erit Dominus rex super omnem terram in die

that day shall there be one LORD, and his name one.

14:10 All the land shall be turned as a plain from Geba illa erit Dominus unus et erit nomen eius unum 14:10 et re-

to Rimmon south of Jerusalem: and it shall be lifted up, and inhabited in her place, from Benjamin’s gate unto the place of the first gate, unto the corner gate, and from the tower of Hananeel unto the king’s winepresses.

uertetur omnis terra usque ad desertum de colle Remmon ad austrum Hierusalem et exaltabitur et habitabit in loco suo a porta Beniamin usque ad locum portae Prioris usque ad portam Angulorum et a turre Ananehel usque ad torcularia

14:11 And men shall dwell in it, and there shall be no more regis 14:11 et habitabunt in ea et anathema non erit amputter destruction; but Jerusalem shall be safely inhabited.

14:12 And this shall be the plague wherewith the lius sed sedebit Hierusalem secura 14:12 et haec erit plaga

LORD will smite all the people that have fought against Jerusalem; Their flesh shall consume away while they stand upon their feet, and their eyes shall consume away in their holes, and their tongue shall consume away in their mouth.

qua percutiet Dominus omnes gentes quae pugnauerunt ad-

Prophetia Zachariae

uersus Hierusalem tabescet caro uniuscuiusque stantis super pedes suos et oculi eius contabescent in foraminibus suis et lingua eorum contabescet in ore suo 14:13 in die illo erit 14:13 And it shall come to pass in that day, that a great

tumult from the LORD shall be among them; and they shall lay hold every one on the hand of his neighbour, and

tumultus Domini magnus in eis et adprehendet uir manum his hand shall rise up against the hand of his neighbour.

proximi sui et conseretur manus eius super manum proximi sui 14:14 sed et Iudas pugnabit aduersus Hierusalem et con- 14:14 And Judah also shall fight at Jerusalem; and the

wealth of all the heathen round about shall be gathered together, gold, and silver, and apparel, in great abundance.

gregabuntur diuitiae omnium gentium in circuitu aurum et argentum et uestes multae satis 14:15 et sic erit ruina equi 14:15 And so shall be the plague of the horse, of the mule,

of the camel, and of the ass, and of all the beasts that shall be in these tents, as this plague.

et muli cameli et asini et omnium iumentorum quae fuerint in castris illis sicut ruina haec 14:16 et omnes qui reliqui 14:16 And it shall come to pass, that every one that is left

of all the nations which came against Jerusalem shall even go up from year to year to worship the King, the LORD of

fuerint de uniuersis gentibus quae uenerint contra Hierusa- hosts, and to keep the feast of tabernacles. lem ascendent ab anno in annum ut adorent regem Dominum exercituum et celebrent festiuitatem tabernaculorum 14:17 14:17 And it shall be, that whoso will not come up of all

the families of the earth unto Jerusalem to worship the King, the LORD of hosts, even upon them shall be no rain.

et erit qui non ascenderit de familiis terrae ad Hierusalem ut adoret regem Dominum exercituum non erit super eos imber 14:18 quod si et familia Aegypti non ascenderit et non uene- 14:18 And if the family of Egypt go not up, and come not,

that have no rain; there shall be the plague, wherewith the LORD will smite the heathen that come not up to keep the

rit nec super eos erit sed erit ruina qua percutiet Dominus feast of tabernacles. omnes gentes quae non ascenderint ad celebrandam festiuitatem tabernaculorum 14:19 hoc erit peccatum Aegypti et 14:19 This shall be the punishment of Egypt, and the pun-

ishment of all nations that come not up to keep the feast of tabernacles.

hoc peccatum omnium gentium quae non ascenderint ad celebrandam festiuitatem tabernaculorum 14:20 in die illo erit 14:20 In that day shall there be upon the bells of the

horses, HOLINESS UNTO THE LORD; and the pots in the LORD’s house shall be like the bowls before the altar.

quod super frenum equi est sanctum Domino et erunt lebetes in domo Domini quasi fialae coram altari 14:21 et erit 14:21 Yea, every pot in Jerusalem and in Judah shall be

holiness unto the LORD of hosts: and all they that sacrifice shall come and take of them, and seethe therein: and in that day there shall be no more the Canaanite in the house of the LORD of hosts.

omnis lebes in Hierusalem et in Iuda sanctificatus Domino exercituum et uenient omnes immolantes et sument ex eis et coquent in eis et non erit mercator ultra in domo Domini exercituum in die illo

Biblia Sacra Vulgata

MCCCXXXVI

Prophetia Malachiae

1:1 onus uerbi Domini ad Israhel in manu Malachi 1:2 di- 1:1 The burden of the word of the LORD to Israel by

Malachi.

1:2 I have loved you, saith the LORD. Yet ye say, Wherein hast thou loved us? Was not Esau Jacob’s brother? saith the LORD: yet I loved Jacob,

lexi uos dicit Dominus et dixistis in quo dilexisti nos nonne frater erat Esau Iacob dicit Dominus et dilexi Iacob 1:3 1:3 And I hated Esau, and laid his mountains and his her-

itage waste for the dragons of the wilderness.

Esau autem odio habui et posui montes eius in solitudinem et hereditatem eius in dracones deserti 1:4 quod si dixerit 1:4 Whereas Edom saith, We are impoverished, but we

will return and build the desolate places; thus saith the LORD of hosts, They shall build, but I will throw down; and they shall call them, The border of wickedness, and, The people against whom the LORD hath indignation for ever.

Idumea destructi sumus sed reuertentes aedificabimus quae deserta sunt haec dicit Dominus exercituum isti aedificabunt et ego destruam et uocabuntur Termini impietatis et Populus cui iratus est Dominus usque in aeternum 1:5 et oculi 1:5 And your eyes shall see, and ye shall say, The LORD

will be magnified from the border of Israel.

uestri uidebunt et uos dicetis magnificetur Dominus super terminum Israhel 1:6 filius honorat patrem et seruus domi- 1:6 A son honoureth his father, and a servant his master: if

then I be a father, where is mine honour? and if I be a master, where is my fear? saith the LORD of hosts unto you, O priests, that despise my name. And ye say, Wherein have we despised thy name?

num suum si ergo pater ego sum ubi est honor meus et si dominus ego sum ubi est timor meus dicit Dominus exercituum ad uos o sacerdotes qui despicitis nomen meum et dixistis in quo despeximus nomen tuum 1:7 offertis super 1:7 Ye offer polluted bread upon mine altar; and ye say,

Wherein have we polluted thee? In that ye say, The table of the LORD is contemptible.

altare meum panem pollutum et dicitis in quo polluimus te in eo quod dicitis mensa Domini despecta est 1:8 si offera- 1:8 And if ye offer the blind for sacrifice, is it not evil?

and if ye offer the lame and sick, is it not evil? offer it now unto thy governor; will he be pleased with thee, or

tis caecum ad immolandum nonne malum est et si offeratis accept thy person? saith the LORD of hosts. claudum et languidum nonne malum est offer illud duci tuo si placuerit ei aut si susceperit faciem tuam dicit Dominus exercituum 1:9 et nunc deprecamini uultum Dei ut misere- 1:9 And now, I pray you, beseech God that he will be gra-

cious unto us: this hath been by your means: will he regard your persons? saith the LORD of hosts.

atur uestri de manu enim uestra factum est hoc si quo modo suscipiat facies uestras dicit Dominus exercituum 1:10 quis 1:10 Who is there even among you that would shut the

doors for nought? neither do ye kindle fire on mine altar for nought. I have no pleasure in you, saith the LORD of

est in uobis qui claudat ostia et incendat altare meum gra- hosts, neither will I accept an offering at your hand. tuito non est mihi uoluntas in uobis dicit Dominus exercituum et munus non suscipiam de manu uestra 1:11 ab ortu 1:11 For from the rising of the sun even unto the going

down of the same my name shall be great among the Gentiles; and in every place incense shall be offered unto my name, and a pure offering: for my name shall be great among the heathen, saith the LORD of hosts.

enim solis usque ad occasum magnum est nomen meum in gentibus et in omni loco sacrificatur et offertur nomini meo oblatio munda quia magnum nomen meum in gentibus dicit Dominus exercituum 1:12 et uos polluistis illud in eo quod 1:12 But ye have profaned it, in that ye say, The table of

the LORD is polluted; and the fruit thereof, even his meat, is contemptible.

dicitis mensa Domini contaminata est et quod superponitur contemptibile est cum igni qui illud deuorat 1:13 et dixistis 1:13 Ye said also, Behold, what a weariness is it! and ye

have snuffed at it, saith the LORD of hosts; and ye brought that which was torn, and the lame, and the sick; thus ye brought an offering: should I accept this of your hand? saith the LORD.

ecce de labore et exsuflastis illud dicit Dominus exercituum

Biblia Sacra Vulgata

et intulistis de rapinis claudum et languidum et intulistis munus numquid suscipiam illud de manu uestra dicit Dominus

1:14 But cursed be the deceiver, which hath in his flock a 1:14 maledictus dolosus qui habet in grege suo masculum et

male, and voweth, and sacrificeth unto the LORD a corrupt thing: for I am a great King, saith the LORD of hosts,

and my name is dreadful among the heathen. uotum faciens immolat debile Domino quia rex magnus ego dicit Dominus exercituum et nomen meum horribile in gentibus

2:1 And now, O ye priests, this commandment is for you. 2:1 et nunc ad uos mandatum hoc o sacerdotes 2:2 si no-2:2 If ye will not hear, and if ye will not lay it to heart, to give glory unto my name, saith the LORD of hosts, I will even send a curse upon you, and I will curse your blessings: yea, I have cursed them already, because ye do not lay it to heart.

lueritis audire et si nolueritis ponere super cor ut detis gloriam nomini meo ait Dominus exercituum mittam in uos egestatem et maledicam benedictionibus uestris et maledicam

2:3 Behold, I will corrupt your seed, and spread dung upon illis quoniam non posuistis super cor 2:3 ecce ego proiciam

your faces, even the dung of your solemn feasts; and one shall take you away with it.

uobis brachium et dispergam super uultum uestrum stercus

2:4 And ye shall know that I have sent this commandment sollemnitatum uestrarum et adsumet uos secum 2:4 et scie-

unto you, that my covenant might be with Levi, saith the LORD of hosts.

tis quia misi ad uos mandatum istud ut esset pactum meum

2:5 My covenant was with him of life and peace; and I cum Leui dicit Dominus exercituum 2:5 pactum meum fuit

gave them to him for the fear wherewith he feared me, and was afraid before my name.

cum eo uitae et pacis et dedi ei timorem et timuit me et a fa-

2:6 The law of truth was in his mouth, and iniquity was not cie nominis mei pauebat 2:6 lex ueritatis fuit in ore eius et

found in his lips: he walked with me in peace and equity, and did turn many away from iniquity.

iniquitas non est inuenta in labiis eius in pace et in aequitate

2:7 For the priest’s lips should keep knowledge, and they ambulauit mecum et multos auertit ab iniquitate 2:7 labia

should seek the law at his mouth: for he is the messenger of the LORD of hosts.

enim sacerdotis custodient scientiam et legem requirent ex

2:8 But ye are departed out of the way; ye have caused ore eius quia angelus Domini exercituum est 2:8 uos autem

many to stumble at the law; ye have corrupted the covenant of Levi, saith the LORD of hosts.

recessistis de uia et scandalizastis plurimos in lege irritum

2:9 Therefore have I also made you contemptible and base fecistis pactum Leui dicit Dominus exercituum 2:9 prop-

before all the people, according as ye have not kept my ways, but have been partial in the law.

ter quod et ego dedi uos contemptibiles et humiles omnibus populis sicut non seruastis uias meas et accepistis faciem in

2:10 Have we not all one father? hath not one God created lege 2:10 numquid non pater unus omnium nostrum numus? why do we deal treacherously every man against his brother, by profaning the covenant of our fathers?

quid non Deus unus creauit nos quare ergo despicit unusquisque nostrum fratrem suum uiolans pactum patrum no-

2:11 Judah hath dealt treacherously, and an abomination strorum 2:11 transgressus est Iuda et abominatio facta est in

is committed in Israel and in Jerusalem; for Judah hath profaned the holiness of the LORD which he loved, and

hath married the daughter of a strange god. Israhel et in Hierusalem quia contaminauit Iudas sanctifica-2:12 The LORD will cut off the man that doeth this, the tionem Domini quam dilexit et habuit filiam dei alieni 2:12

master and the scholar, out of the tabernacles of Jacob, and him that offereth an offering unto the LORD of hosts.

disperdat Dominus uirum qui fecerit hoc magistrum et discipulum de tabernaculis Iacob et offerentem munus Domino

Prophetia Malachiae

exercituum 2:13 et hoc rursum fecistis operiebatis lacrimis 2:13 And this have ye done again, covering the altar of

the LORD with tears, with weeping, and with crying out, insomuch that he regardeth not the offering any more, or

altare Domini fletu et mugitu ita ut ultra non respiciam ad sa- receiveth it with good will at your hand. crificium nec accipiam placabile quid de manu uestra 2:14 2:14 Yet ye say, Wherefore? Because the LORD hath been

witness between thee and the wife of thy youth, against whom thou hast dealt treacherously: yet is she thy com-

et dixistis quam ob causam quia Dominus testificatus est in- panion, and the wife of thy covenant. ter te et uxorem pubertatis tuae quam tu despexisti et haec particeps tua et uxor foederis tui 2:15 nonne unus fecit et 2:15 And did not he make one? Yet had he the residue of

the spirit. And wherefore one? That he might seek a godly seed. Therefore take heed to your spirit, and let none deal

residuum spiritus eius est et quid unus quaerit nisi semen treacherously against the wife of his youth. Dei custodite ergo spiritum uestrum et uxorem adulescentiae tuae noli despicere 2:16 cum odio habueris dimitte dicit 2:16 For the LORD, the God of Israel, saith that he hateth

putting away: for one covereth violence with his garment, saith the LORD of hosts: therefore take heed to your spirit,

Dominus Deus Israhel operiet autem iniquitas uestimentum that ye deal not treacherously. eius dicit Dominus exercituum custodite spiritum uestrum et nolite despicere 2:17 laborare fecistis Dominum in sermo- 2:17 Ye have wearied the LORD with your words. Yet ye

say, Wherein have we wearied him? When ye say, Every one that doeth evil is good in the sight of the LORD, and

nibus uestris et dixistis in quo eum fecimus laborare in eo he delighteth in them; or, Where is the God of judgment? cum diceretis omnis qui facit malum bonus est in conspectu Domini et tales ei placent aut certe ubi est Deus iudicii

3:1 ecce ego mittam angelum meum et praeparabit uiam 3:1 Behold, I will send my messenger, and he shall prepare the way before me: and the LORD, whom ye seek, shall suddenly come to his temple, even the messenger of the covenant, whom ye delight in: behold, he shall come, saith the LORD of hosts.

ante faciem meam et statim ueniet ad templum suum dominator quem uos quaeritis et angelus testamenti quem uos uultis ecce uenit dicit Dominus exercituum 3:2 et quis poterit 3:2 But who may abide the day of his coming? and who

shall stand when he appeareth? for he is like a refiner’s fire, and like fullers’ soap:

cogitare diem aduentus eius et quis stabit ad uidendum eum ipse enim quasi ignis conflans et quasi herba fullonum 3:3 3:3 And he shall sit as a refiner and purifier of silver: and

he shall purify the sons of Levi, and purge them as gold and silver, that they may offer unto the LORD an offering

et sedebit conflans et emundans argentum et purgabit filios in righteousness. Leui et colabit eos quasi aurum et quasi argentum et erunt Domino offerentes sacrificia in iustitia 3:4 et placebit Do- 3:4 Then shall the offering of Judah and Jerusalem be

pleasant unto the LORD, as in the days of old, and as in former years.

mino sacrificium Iuda et Hierusalem sicut dies saeculi et sicut anni antiqui 3:5 et accedam ad uos in iudicio et ero testis 3:5 And I will come near to you to judgment; and I will

be a swift witness against the sorcerers, and against the adulterers, and against false swearers, and against those that oppress the hireling in his wages, the widow, and the fatherless, and that turn aside the stranger from his right, and fear not me, saith the LORD of hosts.

uelox maleficis et adulteris et periuris et qui calumniantur mercedem mercennarii uiduas et pupillos et opprimunt peregrinum nec timuerunt me dicit Dominus exercituum 3:6 3:6 For I am the LORD, I change not; therefore ye sons of

Jacob are not consumed.

ego enim Dominus et non mutor et uos filii Iacob non estis consumpti 3:7 a diebus enim patrum uestrorum recessistis 3:7 Even from the days of your fathers ye are gone away

from mine ordinances, and have not kept them. Return unto me, and I will return unto you, saith the LORD of

a legitimis meis et non custodistis reuertimini ad me et re- hosts. But ye said, Wherein shall we return? uertar ad uos dicit Dominus exercituum et dixistis in quo re-

Biblia Sacra Vulgata

3:8 Will a man rob God? Yet ye have robbed me. But ye uertemur 3:8 si adfiget homo Deum quia uos configitis me

say, Wherein have we robbed thee? In tithes and offerings.

et dixistis in quo confiximus te in decimis et in primitiuis

3:9 Ye are cursed with a curse: for ye have robbed me, 3:9 et in penuria uos maledicti estis et me uos configitis gens

even this whole nation. 3:10 Bring ye all the tithes into the storehouse, that there tota 3:10 inferte omnem decimam in horreum et sit cibus may be meat in mine house, and prove me now herewith, saith the LORD of hosts, if I will not open you the windows of heaven, and pour you out a blessing, that there shall not be room enough to receive it.

in domo mea et probate me super hoc dicit Dominus si non aperuero uobis cataractas caeli et effudero uobis benedic-

3:11 And I will rebuke the devourer for your sakes, and tionem usque ad abundantiam 3:11 et increpabo pro uobis

he shall not destroy the fruits of your ground; neither shall your vine cast her fruit before the time in the field, saith

the LORD of hosts. deuorantem et non corrumpet fructum terrae uestrae nec erit 3:12 And all nations shall call you blessed: for ye shall be sterilis uinea in agro dicit Dominus exercituum 3:12 et be-

a delightsome land, saith the LORD of hosts.

atos uos dicent omnes gentes eritis enim uos terra desidera-

3:13 Your words have been stout against me, saith the bilis dicit Dominus exercituum 3:13 inualuerunt super me

LORD. Yet ye say, What have we spoken so much against thee?

uerba uestra dicit Dominus 3:14 et dixistis quid locuti su-

3:14 Ye have said, It is vain to serve God: and what profit is it that we have kept his ordinance, and that we have walked mournfully before the LORD of hosts?

mus contra te dixistis uanus est qui seruit Deo et quod emolumentum quia custodiuimus praecepta eius et quia ambu-

3:15 And now we call the proud happy; yea, they that work lauimus tristes coram Domino exercituum 3:15 ergo nunc

wickedness are set up; yea, they that tempt God are even delivered.

beatos dicimus arrogantes siquidem aedificati sunt facientes

3:16 Then they that feared the LORD spake often one to impietatem et temptauerunt Deum et salui facti sunt 3:16

another: and the LORD hearkened, and heard it, and a book of remembrance was written before him for them

that feared the LORD, and that thought upon his name. tunc locuti sunt timentes Deum unusquisque cum proximo suo et adtendit Dominus et audiuit et scriptus est liber monumenti coram eo timentibus Dominum et cogitantibus nomen

3:17 And they shall be mine, saith the LORD of hosts, in eius 3:17 et erunt mihi ait Dominus exercituum in die qua

that day when I make up my jewels; and I will spare them, as a man spareth his own son that serveth him.

ego facio in peculium et parcam eis sicut parcit uir filio suo

3:18 Then shall ye return, and discern between the righ- seruienti sibi 3:18 et conuertemini et uidebitis quid sit inter

teous and the wicked, between him that serveth God and him that serveth him not.

iustum et impium et inter seruientem Deo et non seruientem ei

4:1 For, behold, the day cometh, that shall burn as an oven; 4:1 ecce enim dies ueniet succensa quasi caminus et erunt and all the proud, yea, and all that do wickedly, shall be stubble: and the day that cometh shall burn them up, saith the LORD of hosts, that it shall leave them neither root nor branch.

omnes superbi et omnes facientes impietatem stipula et inflammabit eos dies ueniens dicit Dominus exercituum quae

4:2 But unto you that fear my name shall the Sun of righ- non relinquet eis radicem et germen 4:2 et orietur uobis ti-

teousness arise with healing in his wings; and ye shall go forth, and grow up as calves of the stall.

mentibus nomen meum sol iustitiae et sanitas in pinnis eius

4:3 And ye shall tread down the wicked; for they shall be et egrediemini et salietis sicut uituli de armento 4:3 et cal-

ashes under the soles of your feet in the day that I shall do this, saith the LORD of hosts.

cabitis impios cum fuerint cinis sub planta pedum uestrorum

4:4 Remember ye the law of Moses my servant, which in die qua ego facio dicit Dominus exercituum 4:4 memen-

I commanded unto him in Horeb for all Israel, with the statutes and judgments.

Prophetia Malachiae

tote legis Mosi serui mei quam mandaui ei in Choreb ad omnem Israhel praecepta et iudicia 4:5 ecce ego mittam uobis 4:5 Behold, I will send you Elijah the prophet before the

coming of the great and dreadful day of the LORD:

Heliam prophetam antequam ueniat dies Domini magnus et horribilis 4:6 et conuertet cor patrum ad filios et cor filio- 4:6 And he shall turn the heart of the fathers to the chil-

dren, and the heart of the children to their fathers, lest I come and smite the earth with a curse.

rum ad patres eorum ne forte ueniam et percutiam terram anathemate

Biblia Sacra Vulgata

MCCCXLII

Euangelium Secundum Matthaeum

1:1 liber generationis Iesu Christi filii Dauid filii Abra- 1:1 The book of the generation of Jesus Christ, the son of

David, the son of Abraham.

ham 1:2 Abraham genuit Isaac Isaac autem genuit Iacob 1:2 Abraham begat Isaac; and Isaac begat Jacob; and Ja-

cob begat Judas and his brethren;

Iacob autem genuit Iudam et fratres eius 1:3 Iudas autem 1:3 And Judas begat Phares and Zara of Thamar; and

Phares begat Esrom; and Esrom begat Aram;

genuit Phares et Zara de Thamar Phares autem genuit Esrom Esrom autem genuit Aram 1:4 Aram autem genuit Amina- 1:4 And Aram begat Aminadab; and Aminadab begat

Naasson; and Naasson begat Salmon;

dab Aminadab autem genuit Naasson Naasson autem genuit Salmon 1:5 Salmon autem genuit Booz de Rachab Booz 1:5 And Salmon begat Booz of Rachab; and Booz begat

Obed of Ruth; and Obed begat Jesse;

autem genuit Obed ex Ruth Obed autem genuit Iesse Iesse autem genuit Dauid regem 1:6 Dauid autem rex genuit Sa- 1:6 And Jesse begat David the king; and David the king

begat Solomon of her that had been the wife of Urias;

lomonem ex ea quae fuit Vriae 1:7 Salomon autem genuit 1:7 And Solomon begat Roboam; and Roboam begat

Abia; and Abia begat Asa;

Roboam Roboam autem genuit Abiam Abia autem genuit Asa 1:8 Asa autem genuit Iosaphat Iosaphat autem genuit 1:8 And Asa begat Josaphat; and Josaphat begat Joram;

and Joram begat Ozias;

Ioram Ioram autem genuit Oziam 1:9 Ozias autem genuit 1:9 And Ozias begat Joatham; and Joatham begat Achaz;

and Achaz begat Ezekias;

Ioatham Ioatham autem genuit Achaz Achaz autem genuit Ezechiam 1:10 Ezechias autem genuit Manassen Manasses 1:10 And Ezekias begat Manasses; and Manasses begat

Amon; and Amon begat Josias;

autem genuit Amon Amon autem genuit Iosiam 1:11 Io- 1:11 And Josias begat Jechonias and his brethren, about

the time they were carried away to Babylon:

sias autem genuit Iechoniam et fratres eius in transmigratione Babylonis 1:12 et post transmigrationem Babylonis 1:12 And after they were brought to Babylon, Jechonias

begat Salathiel; and Salathiel begat Zorobabel;

Iechonias genuit Salathihel Salathihel autem genuit Zorobabel 1:13 Zorobabel autem genuit Abiud Abiud autem ge- 1:13 And Zorobabel begat Abiud; and Abiud begat Eli-

akim; and Eliakim begat Azor;

nuit Eliachim Eliachim autem genuit Azor 1:14 Azor au- 1:14 And Azor begat Sadoc; and Sadoc begat Achim; and

Achim begat Eliud;

tem genuit Saddoc Saddoc autem genuit Achim Achim autem genuit Eliud 1:15 Eliud autem genuit Eleazar Eleazar 1:15 And Eliud begat Eleazar; and Eleazar begat Matthan;

and Matthan begat Jacob;

autem genuit Matthan Matthan autem genuit Iacob 1:16 Ia- 1:16 And Jacob begat Joseph the husband of Mary, of

whom was born Jesus, who is called Christ.

cob autem genuit Ioseph uirum Mariae de qua natus est Iesus qui uocatur Christus 1:17 omnes ergo generationes ab Ab- 1:17 So all the generations from Abraham to David are

fourteen generations; and from David until the carrying away into Babylon are fourteen generations; and from the carrying away into Babylon unto Christ are fourteen generations.

raham usque ad Dauid generationes quattuordecim et a Dauid usque ad transmigrationem Babylonis generationes quattuordecim et a transmigratione Babylonis usque ad Christum generationes quattuordecim 1:18 Christi autem gene- 1:18 Now the birth of Jesus Christ was on this wise:

When as his mother Mary was espoused to Joseph, before they came together, she was found with child of the

ratio sic erat cum esset desponsata mater eius Maria Ioseph Holy Ghost. antequam conuenirent inuenta est in utero habens de Spiritu Sancto 1:19 Ioseph autem uir eius cum esset iustus et 1:19 Then Joseph her husband, being a just man, and not

willing to make her a publick example, was minded to put her away privily.

1:20 But while he thought on these things, behold, the an- nollet eam traducere uoluit occulte dimittere eam 1:20 haec

gel of the LORD appeared unto him in a dream, saying, Joseph, thou son of David, fear not to take unto thee Mary thy wife: for that which is conceived in her is of the Holy Ghost.

autem eo cogitante ecce angelus Domini in somnis apparuit ei dicens Ioseph fili Dauid noli timere accipere Mariam coniugem tuam quod enim in ea natum est de Spiritu Sancto est

1:21 And she shall bring forth a son, and thou shalt call his 1:21 pariet autem filium et uocabis nomen eius Iesum ipse name JESUS: for he shall save his people from their sins.

1:22 Now all this was done, that it might be fulfilled which enim saluum faciet populum suum a peccatis eorum 1:22 was spoken of the Lord by the prophet, saying,

hoc autem totum factum est ut adimpleretur id quod dictum

1:23 Behold, a virgin shall be with child, and shall bring est a Domino per prophetam dicentem 1:23 ecce uirgo in forth a son, and they shall call his name Emmanuel, which being interpreted is, God with us.

utero habebit et pariet filium et uocabunt nomen eius Emma-

1:24 Then Joseph being raised from sleep did as the angel nuhel quod est interpretatum Nobiscum Deus 1:24 exsurof the Lord had bidden him, and took unto him his wife:

gens autem Ioseph a somno fecit sicut praecepit ei angelus

1:25 And knew her not till she had brought forth her first- Domini et accepit coniugem suam 1:25 et non cognosceborn son: and he called his name JESUS.

bat eam donec peperit filium suum primogenitum et uocauit nomen eius Iesum

2:1 Now when Jesus was born in Bethlehem of Judaea in 2:1 cum ergo natus esset Iesus in Bethleem Iudaeae in die-

the days of Herod the king, behold, there came wise men from the east to Jerusalem,

bus Herodis regis ecce magi ab oriente uenerunt Hierosoly-

2:2 Saying, Where is he that is born King of the Jews? for mam 2:2 dicentes ubi est qui natus est rex Iudaeorum uiwe have seen his star in the east, and are come to worship him.

dimus enim stellam eius in oriente et uenimus adorare eum

2:3 When Herod the king had heard these things, he was 2:3 audiens autem Herodes rex turbatus est et omnis Hiero-

troubled, and all Jerusalem with him.

2:4 And when he had gathered all the chief priests and solyma cum illo 2:4 et congregans omnes principes sacer-

scribes of the people together, he demanded of them where Christ should be born.

dotum et scribas populi sciscitabatur ab eis ubi Christus na-

2:5 And they said unto him, In Bethlehem of Judaea: for sceretur 2:5 at illi dixerunt ei in Bethleem Iudaeae sic enim

thus it is written by the prophet,

2:6 And thou Bethlehem, in the land of Juda, art not the scriptum est per prophetam 2:6 et tu Bethleem terra Iuda

least among the princes of Juda: for out of thee shall come a Governor, that shall rule my people Israel.

nequaquam minima es in principibus Iuda ex te enim ex-

2:7 Then Herod, when he had privily called the wise men, iet dux qui reget populum meum Israhel 2:7 tunc Herodes enquired of them diligently what time the star appeared.

clam uocatis magis diligenter didicit ab eis tempus stellae

2:8 And he sent them to Bethlehem, and said, Go and quae apparuit eis 2:8 et mittens illos in Bethleem dixit ite et

search diligently for the young child; and when ye have found him, bring me word again, that I may come and

worship him also. interrogate diligenter de puero et cum inueneritis renuntiate

2:9 When they had heard the king, they departed; and, lo, mihi ut et ego ueniens adorem eum 2:9 qui cum audissent the star, which they saw in the east, went before them, till it came and stood over where the young child was.

regem abierunt et ecce stella quam uiderant in oriente antecedebat eos usque dum ueniens staret supra ubi erat puer

2:10 When they saw the star, they rejoiced with exceeding 2:10 uidentes autem stellam gauisi sunt gaudio magno ualde

great joy.

2:11 And when they were come into the house, they saw 2:11 et intrantes domum inuenerunt puerum cum Maria matre

the young child with Mary his mother, and fell down, and worshipped him: and when they had opened their treasures, they presented unto him gifts; gold, and frankincense and myrrh.

Euangelium Secundum Matthaeum

eius et procidentes adorauerunt eum et apertis thesauris suis obtulerunt ei munera aurum tus et murram 2:12 et responso 2:12 And being warned of God in a dream that they should

not return to Herod, they departed into their own country another way.

accepto in somnis ne rediient ad Herodem per aliam uiam reuersi sunt in regionem suam 2:13 qui cum recessissent 2:13 And when they were departed, behold, the angel of

the Lord appeareth to Joseph in a dream, saying, Arise, and take the young child and his mother, and flee into Egypt, and be thou there until I bring thee word: for Herod will seek the young child to destroy him.

ecce angelus Domini apparuit in somnis Ioseph dicens surge et accipe puerum et matrem eius et fuge in Aegyptum et esto ibi usque dum dicam tibi futurum est enim ut Herodes quaerat puerum ad perdendum eum 2:14 qui consurgens accepit 2:14 When he arose, he took the young child and his

mother by night, and departed into Egypt:

puerum et matrem eius nocte et recessit in Aegyptum 2:15 2:15 And was there until the death of Herod: that it might

be fulfilled which was spoken of the Lord by the prophet, saying, Out of Egypt have I called my son.

et erat ibi usque ad obitum Herodis ut adimpleretur quod dictum est a Domino per prophetam dicentem ex Aegypto uocaui filium meum 2:16 tunc Herodes uidens quoniam in- 2:16 Then Herod, when he saw that he was mocked of

the wise men, was exceeding wroth, and sent forth, and slew all the children that were in Bethlehem, and in all the coasts thereof, from two years old and under, according to the time which he had diligently enquired of the wise men.

lusus esset a magis iratus est ualde et mittens occidit omnes pueros qui erant in Bethleem et in omnibus finibus eius a bimatu et infra secundum tempus quod exquisierat a magis 2:17 tunc adimpletum est quod dictum est per Hieremiam 2:17 Then was fulfilled that which was spoken by Jeremy

the prophet, saying,

prophetam dicentem 2:18 uox in Rama audita est ploratus 2:18 In Rama was there a voice heard, lamentation, and

weeping, and great mourning, Rachel weeping for her children, and would not be comforted, because they are

et ululatus multus Rachel plorans filios suos et noluit con- not. solari quia non sunt 2:19 defuncto autem Herode ecce ap- 2:19 But when Herod was dead, behold, an angel of the

Lord appeareth in a dream to Joseph in Egypt,

paruit angelus Domini in somnis Ioseph in Aegypto 2:20 2:20 Saying, Arise, and take the young child and his

mother, and go into the land of Israel: for they are dead which sought the young child’s life.

dicens surge et accipe puerum et matrem eius et uade in terram Israhel defuncti sunt enim qui quaerebant animam pueri 2:21 qui surgens accepit puerum et matrem eius et uenit in 2:21 And he arose, and took the young child and his

mother, and came into the land of Israel.

terram Israhel 2:22 audiens autem quod Archelaus regnaret 2:22 But when he heard that Archelaus did reign in Ju-

daea in the room of his father Herod, he was afraid to go thither: notwithstanding, being warned of God in a dream,

in Iudaea pro Herode patre suo timuit illo ire et admonitus he turned aside into the parts of Galilee: in somnis secessit in partes Galilaeae 2:23 et ueniens habi- 2:23 And he came and dwelt in a city called Nazareth: that

it might be fulfilled which was spoken by the prophets, He shall be called a Nazarene.

tauit in ciuitate quae uocatur Nazareth ut adimpleretur quod dictum est per prophetas quoniam Nazareus uocabitur

3:1 in diebus autem illis uenit Iohannes Baptista praedi- 3:1 In those days came John the Baptist, preaching in the

wilderness of Judaea,

cans in deserto Iudaeae 3:2 et dicens paenitentiam agite ad- 3:2 And saying, Repent ye: for the kingdom of heaven is

at hand.

propinquauit enim regnum caelorum 3:3 hic est enim qui 3:3 For this is he that was spoken of by the prophet Esaias, saying, The voice of one crying in the wilderness, Prepare ye the way of the Lord, make his paths straight.

dictus est per Esaiam prophetam dicentem uox clamantis in deserto parate uiam Domini rectas facite semitas eius 3:4 3:4 And the same John had his raiment of camel’s hair, and

a leathern girdle about his loins; and his meat was locusts and wild honey.

Biblia Sacra Vulgata

ipse autem Iohannes habebat uestimentum de pilis camelorum et zonam pelliciam circa lumbos suos esca autem eius

3:5 Then went out to him Jerusalem, and all Judaea, and erat lucustae et mel siluestre 3:5 tunc exiebat ad eum Hie-

all the region round about Jordan,

3:6 And were baptized of him in Jordan, confessing their rosolyma et omnis Iudaea et omnis regio circa Iordanen 3:6

sins.

et baptizabantur in Iordane ab eo confitentes peccata sua

3:7 But when he saw many of the Pharisees and Sadducees 3:7 uidens autem multos Pharisaeorum et Sadducaeorum ue-

come to his baptism, he said unto them, O generation of vipers, who hath warned you to flee from the wrath to

come? nientes ad baptismum suum dixit eis progenies uiperarum 3:8 Bring forth therefore fruits meet for repentance: quis demonstrauit uobis fugere a futura ira 3:8 facite ergo 3:9 And think not to say within yourselves, We have Abra- fructum dignum paenitentiae 3:9 et ne uelitis dicere intra

ham to our father: for I say unto you, that God is able of these stones to raise up children unto Abraham.

uos patrem habemus Abraham dico enim uobis quoniam po-

3:10 And now also the axe is laid unto the root of the trees: test Deus de lapidibus istis suscitare filios Abrahae 3:10

therefore every tree which bringeth not forth good fruit is hewn down, and cast into the fire.

iam enim securis ad radicem arborum posita est omnis ergo arbor quae non facit fructum bonum exciditur et in ignem

3:11 I indeed baptize you with water unto repentance. but mittitur 3:11 ego quidem uos baptizo in aqua in paenitenhe that cometh after me is mightier than I, whose shoes I am not worthy to bear: he shall baptize you with the Holy

Ghost, and with fire: tiam qui autem post me uenturus est fortior me est cuius non

sum dignus calciamenta portare ipse uos baptizabit in Spiritu

3:12 Whose fan is in his hand, and he will throughly purge Sancto et igni 3:12 cuius uentilabrum in manu sua et perhis floor, and gather his wheat into the garner; but he will burn up the chaff with unquenchable fire.

mundabit aream suam et congregabit triticum suum in hor-

3:13 Then cometh Jesus from Galilee to Jordan unto John, reum paleas autem conburet igni inextinguibili 3:13 tunc

to be baptized of him.

uenit Iesus a Galilaea in Iordanen ad Iohannem ut baptizare-

3:14 But John forbad him, saying, I have need to be bap- tur ab eo 3:14 Iohannes autem prohibebat eum dicens ego a

tized of thee, and comest thou to me?

3:15 And Jesus answering said unto him, Suffer it to be te debeo baptizari et tu uenis ad me 3:15 respondens autem

so now: for thus it becometh us to fulfil all righteousness. Then he suffered him.

Iesus dixit ei sine modo sic enim decet nos implere omnem

3:16 And Jesus, when he was baptized, went up straight- iustitiam tunc dimisit eum 3:16 baptizatus autem confestim

way out of the water: and, lo, the heavens were opened unto him, and he saw the Spirit of God descending like a

dove, and lighting upon him: ascendit de aqua et ecce aperti sunt ei caeli et uidit Spiri-

tum Dei descendentem sicut columbam uenientem super se

3:17 And lo a voice from heaven, saying, This is my 3:17 et ecce uox de caelis dicens hic est Filius meus dilectus

beloved Son, in whom I am well pleased.

in quo mihi conplacui

4:1 Then was Jesus led up of the spirit into the wilderness 4:1 tunc Iesus ductus est in desertum ab Spiritu ut tempto be tempted of the devil.

4:2 And when he had fasted forty days and forty nights, he taretur a diabolo 4:2 et cum ieiunasset quadraginta diebus

was afterward an hungred.

4:3 And when the tempter came to him, he said, If thou et quadraginta noctibus postea esuriit 4:3 et accedens tem-

be the Son of God, command that these stones be made bread.

ptator dixit ei si Filius Dei es dic ut lapides isti panes fiant

4:4 But he answered and said, It is written, Man shall not 4:4 qui respondens dixit scriptum est non in pane solo ui-

live by bread alone, but by every word that proceedeth out of the mouth of God.

Euangelium Secundum Matthaeum

uet homo sed in omni uerbo quod procedit de ore Dei 4:5 4:5 Then the devil taketh him up into the holy city, and

setteth him on a pinnacle of the temple,

tunc adsumit eum diabolus in sanctam ciuitatem et statuit eum supra pinnaculum templi 4:6 et dixit ei si Filius Dei es 4:6 And saith unto him, If thou be the Son of God, cast

thyself down: for it is written, He shall give his angels charge concerning thee: and in their hands they shall bear

mitte te deorsum scriptum est enim quia angelis suis man- thee up, lest at any time thou dash thy foot against a stone.

dabit de te et in manibus tollent te ne forte offendas ad la-

pidem pedem tuum 4:7 ait illi Iesus rursum scriptum est 4:7 Jesus said unto him, It is written again, Thou shalt not

tempt the Lord thy God.

non temptabis Dominum Deum tuum 4:8 iterum adsumit 4:8 Again, the devil taketh him up into an exceeding high mountain, and sheweth him all the kingdoms of the world, and the glory of them;

eum diabolus in montem excelsum ualde et ostendit ei omnia regna mundi et gloriam eorum 4:9 et dixit illi haec tibi 4:9 And saith unto him, All these things will I give thee, if

thou wilt fall down and worship me.

omnia dabo si cadens adoraueris me 4:10 tunc dicit ei Ie- 4:10 Then saith Jesus unto him, Get thee hence, Satan: for it is written, Thou shalt worship the Lord thy God, and him only shalt thou serve.

sus uade Satanas scriptum est Dominum Deum tuum adorabis et illi soli seruies 4:11 tunc reliquit eum diabolus et 4:11 Then the devil leaveth him, and, behold, angels came

and ministered unto him.

ecce angeli accesserunt et ministrabant ei 4:12 cum autem 4:12 Now when Jesus had heard that John was cast into

prison, he departed into Galilee;

audisset quod Iohannes traditus esset secessit in Galilaeam 4:13 et relicta ciuitate Nazareth uenit et habitauit in Caphar- 4:13 And leaving Nazareth, he came and dwelt in Caper-

naum, which is upon the sea coast, in the borders of Zabulon and Nephthalim:

naum maritimam in finibus Zabulon et Nepthalim 4:14 ut

4:14 That it might be fulfilled which was spoken by Esaias the prophet, saying,

adimpleretur quod dictum est per Esaiam prophetam 4:15 _{4:15 The land of Zabulon, and the land of Nephthalim, by}

the way of the sea, beyond Jordan, Galilee of the Gentiles;

terra Zabulon et terra Nepthalim uia maris trans Iordanen Galilaeae gentium 4:16 populus qui sedebat in tenebris lu- 4:16 The people which sat in darkness saw great light; and

to them which sat in the region and shadow of death light is sprung up.

cem uidit magnam et sedentibus in regione et umbra mortis lux orta est eis 4:17 exinde coepit Iesus praedicare et di- 4:17 From that time Jesus began to preach, and to say,

Repent: for the kingdom of heaven is at hand.

cere paenitentiam agite adpropinquauit enim regnum caelorum 4:18 ambulans autem iuxta mare Galilaeae uidit duos 4:18 And Jesus, walking by the sea of Galilee, saw two

brethren, Simon called Peter, and Andrew his brother, casting a net into the sea: for they were fishers.

fratres Simonem qui uocatur Petrus et Andream fratrem eius mittentes rete in mare erant enim piscatores 4:19 et ait illis 4:19 And he saith unto them, Follow me, and I will make

you fishers of men.

uenite post me et faciam uos fieri piscatores hominum 4:20 4:20 And they straightway left their nets, and followed

him.

at illi continuo relictis retibus secuti sunt eum 4:21 et pro- 4:21 And going on from thence, he saw other two brethren,

James the son of Zebedee, and John his brother, in a ship with Zebedee their father, mending their nets; and

cedens inde uidit alios duos fratres Iacobum Zebedaei et Io- he called them. hannem fratrem eius in naui cum Zebedaeo patre eorum reficientes retia sua et uocauit eos 4:22 illi autem statim relic- 4:22 And they immediately left the ship and their father,

and followed him.

tis retibus et patre secuti sunt eum 4:23 et circumibat Iesus 4:23 And Jesus went about all Galilee, teaching in their

synagogues, and preaching the gospel of the kingdom, and healing all manner of sickness and all manner of disease

totam Galilaeam docens in synagogis eorum et praedicans among the people. euangelium regni et sanans omnem languorem et omnem in-

Biblia Sacra Vulgata

4:24 And his fame went throughout all Syria: and they firmitatem in populo 4:24 et abiit opinio eius in totam Sy-

brought unto him all sick people that were taken with divers diseases and torments, and those which were possessed with devils, and those which were lunatick, and those that had the palsy; and he healed them.

riam et obtulerunt ei omnes male habentes uariis languoribus et tormentis conprehensos et qui daemonia habebant et

4:25 And there followed him great multitudes of people lunaticos et paralyticos et curauit eos 4:25 et secutae sunt

from Galilee, and from Decapolis, and from Jerusalem, and from Judaea, and from beyond Jordan.

eum turbae multae de Galilaea et Decapoli et Hierosolymis et Iudaea et de trans Iordanen

5:1 And seeing the multitudes, he went up into a mountain: 5:1 uidens autem turbas ascendit in montem et cum seand when he was set, his disciples came unto him:

5:2 And he opened his mouth, and taught them, saying, disset accesserunt ad eum discipuli eius 5:2 et aperiens os

5:3 Blessed are the poor in spirit: for theirs is the kingdom suum docebat eos dicens 5:3 beati pauperes spiritu quoof heaven. 5:4 Blessed are they that mourn: for they shall be com- niam ipsorum est regnum caelorum 5:4 beati mites quoforted. 5:5 Blessed are the meek: for they shall inherit the earth. niam ipsi possidebunt terram 5:5 beati qui lugent quoniam

5:6 Blessed are they which do hunger and thirst after righ- ipsi consolabuntur 5:6 beati qui esuriunt et sitiunt iustitiam

teousness: for they shall be filled.

5:7 Blessed are the merciful: for they shall obtain mercy. quoniam ipsi saturabuntur 5:7 beati misericordes quia ipsi

5:8 Blessed are the pure in heart: for they shall see God. misericordiam consequentur 5:8 beati mundo corde quo-5:9 Blessed are the peacemakers: for they shall be called niam ipsi Deum uidebunt 5:9 beati pacifici quoniam filii the children of God.

5:10 Blessed are they which are persecuted for righteous- Dei uocabuntur 5:10 beati qui persecutionem patiuntur pro-

ness’ sake: for theirs is the kingdom of heaven.

5:11 Blessed are ye, when men shall revile you, and per- pter iustitiam quoniam ipsorum est regnum caelorum 5:11

secute you, and shall say all manner of evil against you falsely, for my sake.

beati estis cum maledixerint uobis et persecuti uos fuerint et dixerint omne malum aduersum uos mentientes propter me

5:12 Rejoice, and be exceeding glad: for great is your re- 5:12 gaudete et exultate quoniam merces uestra copiosa est in

ward in heaven: for so persecuted they the prophets which were before you.

caelis sic enim persecuti sunt prophetas qui fuerunt ante uos

5:13 Ye are the salt of the earth: but if the salt have lost his 5:13 uos estis sal terrae quod si sal euanuerit in quo sallie-

savour, wherewith shall it be salted? it is thenceforth good for nothing, but to be cast out, and to be trodden under foot

of men. tur ad nihilum ualet ultra nisi ut mittatur foras et conculcetur 5:14 Ye are the light of the world. A city that is set on an ab hominibus 5:14 uos estis lux mundi non potest ciuitas

hill cannot be hid.

5:15 Neither do men light a candle, and put it under a abscondi supra montem posita 5:15 neque accendunt lucer-

bushel, but on a candlestick; and it giveth light unto all that are in the house.

nam et ponunt eam sub modio sed super candelabrum ut lu-

5:16 Let your light so shine before men, that they may ceat omnibus qui in domo sunt 5:16 sic luceat lux uestra

see your good works, and glorify your Father which is in heaven.

coram hominibus ut uideant uestra bona opera et glorificent

5:17 Think not that I am come to destroy the law, or the Patrem uestrum qui in caelis est 5:17 nolite putare quoniam

prophets: I am not come to destroy, but to fulfil.

ueni soluere legem aut prophetas non ueni soluere sed adim-

5:18 For verily I say unto you, Till heaven and earth pass, plere 5:18 amen quippe dico uobis donec transeat caelum one jot or one tittle shall in no wise pass from the law, till all be fulfilled.

et terra iota unum aut unus apex non praeteribit a lege do-

5:19 Whosoever therefore shall break one of these least nec omnia fiant 5:19 qui ergo soluerit unum de mandatis commandments, and shall teach men so, he shall be called the least in the kingdom of heaven: but whosoever shall do and teach them, the same shall be called great in the kingdom of heaven.

Euangelium Secundum Matthaeum

istis minimis et docuerit sic homines minimus uocabitur in regno caelorum qui autem fecerit et docuerit hic magnus uocabitur in regno caelorum 5:20 dico enim uobis quia nisi 5:20 For I say unto you, That except your righteousness

shall exceed the righteousness of the scribes and Pharisees, ye shall in no case enter into the kingdom of heaven.

abundauerit iustitia uestra plus quam scribarum et Pharisaeorum non intrabitis in regnum caelorum 5:21 audistis quia 5:21 Ye have heard that it was said of them of old time,

Thou shalt not kill; and whosoever shall kill shall be in danger of the judgment:

dictum est antiquis non occides qui autem occiderit reus erit iudicio 5:22 ego autem dico uobis quia omnis qui irascitur 5:22 But I say unto you, That whosoever is angry with his

brother without a cause shall be in danger of the judgment: and whosoever shall say to his brother, Raca, shall be in danger of the council: but whosoever shall say, Thou fool, shall be in danger of hell fire.

fratri suo reus erit iudicio qui autem dixerit fratri suo racha reus erit concilio qui autem dixerit fatue reus erit gehennae

ignis 5:23 si ergo offeres munus tuum ad altare et ibi re- 5:23 Therefore if thou bring thy gift to the altar, and there

rememberest that thy brother hath ought against thee;

cordatus fueris quia frater tuus habet aliquid aduersum te 5:24 relinque ibi munus tuum ante altare et uade prius recon- 5:24 Leave there thy gift before the altar, and go thy way;

first be reconciled to thy brother, and then come and offer thy gift.

ciliare fratri tuo et tunc ueniens offers munus tuum 5:25

5:25 Agree with thine adversary quickly, whiles thou art in the way with him; lest at any time the adversary deliver thee to the judge, and the judge deliver thee to the officer, and thou be cast into prison.

esto consentiens aduersario tuo cito dum es in uia cum eo ne forte tradat te aduersarius iudici et iudex tradat te ministro et in carcerem mittaris 5:26 amen dico tibi non exies 5:26 Verily I say unto thee, Thou shalt by no means come

out thence, till thou hast paid the uttermost farthing.

inde donec reddas nouissimum quadrantem 5:27 audistis 5:27 Ye have heard that it was said by them of old time,

Thou shalt not commit adultery:

quia dictum est antiquis non moechaberis 5:28 ego autem 5:28 But I say unto you, That whosoever looketh on a

woman to lust after her hath committed adultery with her already in his heart.

dico uobis quoniam omnis qui uiderit mulierem ad concupiscendum eam iam moechatus est eam in corde suo 5:29 5:29 And if thy right eye offend thee, pluck it out, and

cast it from thee: for it is profitable for thee that one of thy members should perish, and not that thy whole body

quod si oculus tuus dexter scandalizat te erue eum et proice should be cast into hell. abs te expedit enim tibi ut pereat unum membrorum tuorum

quam totum corpus tuum mittatur in gehennam 5:30 et si 5:30 And if thy right hand offend thee, cut it off, and cast it from thee: for it is profitable for thee that one of thy members should perish, and not that thy whole body should be

dextera manus tua scandalizat te abscide eam et proice abs cast into hell. te expedit tibi ut pereat unum membrorum tuorum quam to-

tum corpus tuum eat in gehennam 5:31 dictum est autem 5:31 It hath been said, Whosoever shall put away his wife,

let him give her a writing of divorcement:

quicumque dimiserit uxorem suam det illi libellum repudii 5:32 ego autem dico uobis quia omnis qui dimiserit uxorem 5:32 But I say unto you, That whosoever shall put away

his wife, saving for the cause of fornication, causeth her to commit adultery: and whosoever shall marry her that is

suam excepta fornicationis causa facit eam moechari et qui divorced committeth adultery. dimissam duxerit adulterat 5:33 iterum audistis quia dictum 5:33 Again, ye have heard that it hath been said by them

of old time, Thou shalt not forswear thyself, but shalt perform unto the Lord thine oaths:

est antiquis non peierabis reddes autem Domino iuramenta tua 5:34 ego autem dico uobis non iurare omnino neque per 5:34 But I say unto you, Swear not at all; neither by

heaven; for it is God’s throne:

caelum quia thronus Dei est 5:35 neque per terram quia sca- 5:35 Nor by the earth; for it is his footstool: neither by

Jerusalem; for it is the city of the great King.

Biblia Sacra Vulgata

billum est pedum eius neque per Hierosolymam quia ciuitas

5:36 Neither shalt thou swear by thy head, because thou est magni Regis 5:36 neque per caput tuum iuraueris quia

canst not make one hair white or black.

5:37 But let your communication be, Yea, yea; Nay, nay: non potes unum capillum album facere aut nigrum 5:37 sit

for whatsoever is more than these cometh of evil.

autem sermo uester est est non non quod autem his abundan-

5:38 Ye have heard that it hath been said, An eye for an tius est a malo est 5:38 audistis quia dictum est oculum pro

eye, and a tooth for a tooth: 5:39 But I say unto you, That ye resist not evil: but whoso- oculo et dentem pro dente 5:39 ego autem dico uobis non ever shall smite thee on thy right cheek, turn to him the other also.

resistere malo sed si quis te percusserit in dextera maxilla tua

5:40 And if any man will sue thee at the law, and take away praebe illi et alteram 5:40 et ei qui uult tecum iudicio con-

thy coat, let him have thy cloak also.

5:41 And whosoever shall compel thee to go a mile, go tendere et tunicam tuam tollere remitte ei et pallium 5:41

with him twain.

et quicumque te angariauerit mille passus uade cum illo alia

5:42 Give to him that asketh thee, and from him that would duo 5:42 qui petit a te da ei et uolenti mutuari a te ne auer-

borrow of thee turn not thou away.

5:43 Ye have heard that it hath been said, Thou shalt love taris 5:43 audistis quia dictum est diliges proximum tuum

thy neighbour, and hate thine enemy.

5:44 But I say unto you, Love your enemies, bless them et odio habebis inimicum tuum 5:44 ego autem dico uobis

that curse you, do good to them that hate you, and pray for them which despitefully use you, and persecute you;

diligite inimicos uestros benefacite his qui oderunt uos et

5:45 That ye may be the children of your Father which is orate pro persequentibus et calumniantibus uos 5:45 ut sitis

in heaven: for he maketh his sun to rise on the evil and on the good, and sendeth rain on the just and on the unjust.

filii Patris uestri qui in caelis est qui solem suum oriri facit

5:46 For if ye love them which love you, what reward have super bonos et malos et pluit super iustos et iniustos 5:46

ye? do not even the publicans the same?

si enim diligatis eos qui uos diligunt quam mercedem habe-

5:47 And if ye salute your brethren only, what do ye more bitis nonne et publicani hoc faciunt 5:47 et si salutaueritis

than others? do not even the publicans so?

fratres uestros tantum quid amplius facitis nonne et ethnici

5:48 Be ye therefore perfect, even as your Father which is hoc faciunt 5:48 estote ergo uos perfecti sicut et Pater ues-

in heaven is perfect.

ter caelestis perfectus est

6:1 Take heed that ye do not your alms before men, to be 6:1 adtendite ne iustitiam uestram faciatis coram homi-

seen of them: otherwise ye have no reward of your Father which is in heaven.

nibus ut uideamini ab eis alioquin mercedem non habebi-

6:2 Therefore when thou doest thine alms, do not sound tis apud Patrem uestrum qui in caelis est 6:2 cum ergo fa-

a trumpet before thee, as the hypocrites do in the synagogues and in the streets, that they may have glory of men.

Verily I say unto you, They have their reward. cies elemosynam noli tuba canere ante te sicut hypocritae fa-

ciunt in synagogis et in uicis ut honorificentur ab hominibus

6:3 But when thou doest alms, let not thy left hand know amen dico uobis receperunt mercedem suam 6:3 te autem what thy right hand doeth:

faciente elemosynam nesciat sinistra tua quid faciat dextera

6:4 That thine alms may be in secret: and thy Father which tua 6:4 ut sit elemosyna tua in abscondito et Pater tuus qui

seeth in secret himself shall reward thee openly.

6:5 And when thou prayest, thou shalt not be as the hyp- uidet in abscondito reddet tibi 6:5 et cum oratis non eritis

ocrites are: for they love to pray standing in the synagogues and in the corners of the streets, that they may be seen of men. Verily I say unto you, They have their reward.

sicut hypocritae qui amant in synagogis et in angulis platearum stantes orare ut uideantur ab hominibus amen dico uobis

Euangelium Secundum Matthaeum

receperunt mercedem suam 6:6 tu autem cum orabis intra 6:6 But thou, when thou prayest, enter into thy closet, and when thou hast shut thy door, pray to thy Father which is in secret; and thy Father which seeth in secret shall reward

in cubiculum tuum et cluso ostio tuo ora Patrem tuum in ab- thee openly. scondito et Pater tuus qui uidet in abscondito reddet tibi 6:7 6:7 But when ye pray, use not vain repetitions, as the hea-

then do: for they think that they shall be heard for their much speaking.

orantes autem nolite multum loqui sicut ethnici putant enim quia in multiloquio suo exaudiantur 6:8 nolite ergo adsi- 6:8 Be not ye therefore like unto them: for your Father

knoweth what things ye have need of, before ye ask him.

milari eis scit enim Pater uester quibus opus sit uobis antequam petatis eum 6:9 sic ergo uos orabitis Pater noster qui 6:9 After this manner therefore pray ye: Our Father which

art in heaven, Hallowed be thy name.

in caelis es sanctificetur nomen tuum 6:10 ueniat regnum 6:10 Thy kingdom come, Thy will be done in earth, as it

is in heaven.

tuum fiat uoluntas tua sicut in caelo et in terra 6:11 panem 6:11 Give us this day our daily bread. nostrum supersubstantialem da nobis hodie 6:12 et dimitte 6:12 And forgive us our debts, as we forgive our debtors. nobis debita nostra sicut et nos dimisimus debitoribus nostris 6:13 et ne inducas nos in temptationem sed libera nos a malo 6:13 And lead us not into temptation, but deliver us from

evil: For thine is the kingdom, and the power, and the glory, for ever. Amen.

6:14 si enim dimiseritis hominibus peccata eorum dimittet

6:14 For if ye forgive men their trespasses, your heavenly Father will also forgive you:

et uobis Pater uester caelestis delicta uestra 6:15 si autem _{6:15 But if ye forgive not men their trespasses, neither will}

your Father forgive your trespasses.

non dimiseritis hominibus nec Pater uester dimittet peccata uestra 6:16 cum autem ieiunatis nolite fieri sicut hypocritae 6:16 Moreover when ye fast, be not, as the hypocrites, of

a sad countenance: for they disfigure their faces, that they may appear unto men to fast. Verily I say unto you, They

tristes demoliuntur enim facies suas ut pareant hominibus have their reward. ieiunantes amen dico uobis quia receperunt mercedem suam 6:17 tu autem cum ieiunas ungue caput tuum et faciem tuam 6:17 But thou, when thou fastest, anoint thine head, and

wash thy face;

laua 6:18 ne uidearis hominibus ieiunans sed Patri tuo qui 6:18 That thou appear not unto men to fast, but unto thy

Father which is in secret: and thy Father, which seeth in secret, shall reward thee openly.

est in abscondito et Pater tuus qui uidet in abscondito reddet tibi 6:19 nolite thesaurizare uobis thesauros in terra ubi 6:19 Lay not up for yourselves treasures upon earth, where

moth and rust doth corrupt, and where thieves break through and steal:

erugo et tinea demolitur ubi fures effodiunt et furantur 6:20

6:20 But lay up for yourselves treasures in heaven, where neither moth nor rust doth corrupt, and where thieves do not break through nor steal:

thesaurizate autem uobis thesauros in caelo ubi neque erugo neque tinea demolitur et ubi fures non effodiunt nec furantur 6:21 ubi enim est thesaurus tuus ibi est et cor tuum 6:22 lu- 6:21 For where your treasure is, there will your heart be

also.

6:22 The light of the body is the eye: if therefore thine eye be single, thy whole body shall be full of light.

cerna corporis est oculus si fuerit oculus tuus simplex totum corpus tuum lucidum erit 6:23 si autem oculus tuus nequam 6:23 But if thine eye be evil, thy whole body shall be full

of darkness. If therefore the light that is in thee be darkness, how great is that darkness!

fuerit totum corpus tuum tenebrosum erit si ergo lumen quod in te est tenebrae sunt tenebrae quantae erunt 6:24 nemo 6:24 No man can serve two masters: for either he will hate

the one, and love the other; or else he will hold to the one, and despise the other. Ye cannot serve God and mammon.

potest duobus dominis seruire aut enim unum odio habebit et alterum diliget aut unum sustinebit et alterum contemnet non potestis Deo seruire et mamonae 6:25 ideo dico uobis 6:25 Therefore I say unto you, Take no thought for your

life, what ye shall eat, or what ye shall drink; nor yet for your body, what ye shall put on. Is not the life more than meat, and the body than raiment?

Biblia Sacra Vulgata

ne solliciti sitis animae uestrae quid manducetis neque corpori uestro quid induamini nonne anima plus est quam esca

6:26 Behold the fowls of the air: for they sow not, nei- et corpus plus est quam uestimentum 6:26 respicite uolati-

ther do they reap, nor gather into barns; yet your heavenly Father feedeth them. Are ye not much better than they?

lia caeli quoniam non serunt neque metunt neque congregant in horrea et Pater uester caelestis pascit illa nonne uos ma-

6:27 Which of you by taking thought can add one cubit gis pluris estis illis 6:27 quis autem uestrum cogitans potest

unto his stature?

6:28 And why take ye thought for raiment? Consider the adicere ad staturam suam cubitum unum 6:28 et de uesti-

lilies of the field, how they grow; they toil not, neither do they spin:

mento quid solliciti estis considerate lilia agri quomodo cre-

6:29 And yet I say unto you, That even Solomon in all his scunt non laborant nec nent 6:29 dico autem uobis quoniam

glory was not arrayed like one of these.

nec Salomon in omni gloria sua coopertus est sicut unum ex

6:30 Wherefore, if God so clothe the grass of the field, istis 6:30 si autem faenum agri quod hodie est et cras in

which to day is, and to morrow is cast into the oven, shall he not much more clothe you, O ye of little faith?

clibanum mittitur Deus sic uestit quanto magis uos minimae

6:31 Therefore take no thought, saying, What shall we fidei 6:31 nolite ergo solliciti esse dicentes quid manduca-

eat? or, What shall we drink? or, Wherewithal shall we be clothed?

bimus aut quid bibemus aut quo operiemur 6:32 haec enim

6:32 (For after all these things do the Gentiles seek:) for your heavenly Father knoweth that ye have need of all these things.

omnia gentes inquirunt scit enim Pater uester quia his om-

6:33 But seek ye first the kingdom of God, and his righ- nibus indigetis 6:33 quaerite autem primum regnum et ius-

teousness; and all these things shall be added unto you.

6:34 Take therefore no thought for the morrow: for the titiam eius et omnia haec adicientur uobis 6:34 nolite ergo

morrow shall take thought for the things of itself. Sufficient unto the day is the evil thereof.

esse solliciti in crastinum crastinus enim dies sollicitus erit sibi ipse sufficit diei malitia sua

7:1 Judge not, that ye be not judged. 7:1 nolite iudicare ut non iudicemini 7:2 in quo enim iu-

7:2 For with what judgment ye judge, ye shall be judged: and with what measure ye mete, it shall be measured to you again.

dicio iudicaueritis iudicabimini et in qua mensura mensi fue-

7:3 And why beholdest thou the mote that is in thy ritis metietur uobis 7:3 quid autem uides festucam in oculo

brother’s eye, but considerest not the beam that is in thine own eye?

fratris tui et trabem in oculo tuo non uides 7:4 aut quomodo

7:4 Or how wilt thou say to thy brother, Let me pull out the mote out of thine eye; and, behold, a beam is in thine own eye?

dicis fratri tuo sine eiciam festucam de oculo tuo et ecce tra-

7:5 Thou hypocrite, first cast out the beam out of thine bis est in oculo tuo 7:5 hypocrita eice primum trabem de

own eye; and then shalt thou see clearly to cast out the mote out of thy brother’s eye.

oculo tuo et tunc uidebis eicere festucam de oculo fratris

7:6 Give not that which is holy unto the dogs, neither cast tui 7:6 nolite dare sanctum canibus neque mittatis margari-

ye your pearls before swine, lest they trample them under their feet, and turn again and rend you.

tas uestras ante porcos ne forte conculcent eas pedibus suis

7:7 Ask, and it shall be given you; seek, and ye shall find; et conuersi disrumpant uos 7:7 petite et dabitur uobis quae-

knock, and it shall be opened unto you:

7:8 For every one that asketh receiveth; and he that seeketh rite et inuenietis pulsate et aperietur uobis 7:8 omnis enim

findeth; and to him that knocketh it shall be opened.

qui petit accipit et qui quaerit inuenit et pulsanti aperietur

7:9 Or what man is there of you, whom if his son ask 7:9 aut quis est ex uobis homo quem si petierit filius suus pa-

bread, will he give him a stone?

7:10 Or if he ask a fish, will he give him a serpent? nem numquid lapidem porriget ei 7:10 aut si piscem petet

Euangelium Secundum Matthaeum

numquid serpentem porriget ei 7:11 si ergo uos cum sitis 7:11 If ye then, being evil, know how to give good gifts

unto your children, how much more shall your Father which is in heaven give good things to them that ask him?

mali nostis bona dare filiis uestris quanto magis Pater uester qui in caelis est dabit bona petentibus se 7:12 omnia ergo 7:12 Therefore all things whatsoever ye would that men

should do to you, do ye even so to them: for this is the law and the prophets.

quaecumque uultis ut faciant uobis homines et uos facite eis

haec est enim lex et prophetae 7:13 intrate per angustam 7:13 Enter ye in at the strait gate: for wide is the gate, and broad is the way, that leadeth to destruction, and many there be which go in thereat:

portam quia lata porta et spatiosa uia quae ducit ad perditionem et multi sunt qui intrant per eam 7:14 quam angusta 7:14 Because strait is the gate, and narrow is the way,

which leadeth unto life, and few there be that find it.

porta et arta uia quae ducit ad uitam et pauci sunt qui inueniunt eam 7:15 adtendite a falsis prophetis qui ueniunt ad 7:15 Beware of false prophets, which come to you in

sheep’s clothing, but inwardly they are ravening wolves.

uos in uestimentis ouium intrinsecus autem sunt lupi rapaces 7:16 a fructibus eorum cognoscetis eos numquid colli- 7:16 Ye shall know them by their fruits. Do men gather

grapes of thorns, or figs of thistles?

gunt de spinis uuas aut de tribulis ficus 7:17 sic omnis arbor 7:17 Even so every good tree bringeth forth good fruit; but

a corrupt tree bringeth forth evil fruit.

bona fructus bonos facit mala autem arbor fructus malos fa-

cit 7:18 non potest arbor bona fructus malos facere neque 7:18 A good tree cannot bring forth evil fruit, neither can

a corrupt tree bring forth good fruit.

arbor mala fructus bonos facere 7:19 omnis arbor quae non 7:19 Every tree that bringeth not forth good fruit is hewn

down, and cast into the fire.

facit fructum bonum exciditur et in ignem mittitur 7:20 igi- 7:20 Wherefore by their fruits ye shall know them. tur ex fructibus eorum cognoscetis eos 7:21 non omnis qui 7:21 Not every one that saith unto me, Lord, Lord, shall

enter into the kingdom of heaven; but he that doeth the will of my Father which is in heaven.

dicit mihi Domine Domine intrabit in regnum caelorum sed qui facit uoluntatem Patris mei qui in caelis est ipse intrabit in regnum caelorum 7:22 multi dicent mihi in illa die Do- 7:22 Many will say to me in that day, Lord, Lord, have we

not prophesied in thy name? and in thy name have cast out devils? and in thy name done many wonderful works?

mine Domine nonne in nomine tuo prophetauimus et in tuo nomine daemonia eiecimus et in tuo nomine uirtutes multas fecimus 7:23 et tunc confitebor illis quia numquam noui 7:23 And then will I profess unto them, I never knew you:

depart from me, ye that work iniquity.

uos discedite a me qui operamini iniquitatem 7:24 omnis 7:24 Therefore whosoever heareth these sayings of mine, and doeth them, I will liken him unto a wise man, which built his house upon a rock:

ergo qui audit uerba mea haec et facit ea adsimilabitur uiro sapienti qui aedificauit domum suam supra petram 7:25 et 7:25 And the rain descended, and the floods came, and the

winds blew, and beat upon that house; and it fell not: for it was founded upon a rock.

descendit pluuia et uenerunt flumina et flauerunt uenti et inruerunt in domum illam et non cecidit fundata enim erat super petram 7:26 et omnis qui audit uerba mea haec et non 7:26 And every one that heareth these sayings of mine,

and doeth them not, shall be likened unto a foolish man, which built his house upon the sand:

facit ea similis erit uiro stulto qui aedificauit domum suam supra harenam 7:27 et descendit pluuia et uenerunt flumina 7:27 And the rain descended, and the floods came, and the

winds blew, and beat upon that house; and it fell: and great was the fall of it.

et flauerunt uenti et inruerunt in domum illam et cecidit et fuit ruina eius magna 7:28 et factum est cum consummas- 7:28 And it came to pass, when Jesus had ended these say-

ings, the people were astonished at his doctrine:

Biblia Sacra Vulgata

set Iesus uerba haec admirabantur turbae super doctrinam

7:29 For he taught them as one having authority, and not eius 7:29 erat enim docens eos sicut potestatem habens non

as the scribes.

sicut scribae eorum et Pharisaei

8:1 When he was come down from the mountain, great 8:1 cum autem descendisset de monte secutae sunt eum

multitudes followed him.

8:2 And, behold, there came a leper and worshipped him, turbae multae 8:2 et ecce leprosus ueniens adorabat eum

saying, Lord, if thou wilt, thou canst make me clean. 8:3 And Jesus put forth his hand, and touched him, saying, dicens Domine si uis potes me mundare 8:3 et extendens I will; be thou clean. And immediately his leprosy was cleansed.

manum tetigit eum Iesus dicens uolo mundare et confestim

8:4 And Jesus saith unto him, See thou tell no man; but go mundata est lepra eius 8:4 et ait illi Iesus uide nemini dixe-

thy way, shew thyself to the priest, and offer the gift that Moses commanded, for a testimony unto them.

ris sed uade ostende te sacerdoti et offer munus quod prae-

8:5 And when Jesus was entered into Capernaum, there cepit Moses in testimonium illis 8:5 cum autem introisset

came unto him a centurion, beseeching him, 8:6 And saying, Lord, my servant lieth at home sick of the Capharnaum accessit ad eum centurio rogans eum 8:6 et palsy, grievously tormented.

dicens Domine puer meus iacet in domo paralyticus et male

8:7 And Jesus saith unto him, I will come and heal him. torquetur 8:7 et ait illi Iesus ego ueniam et curabo eum 8:8

8:8 The centurion answered and said, Lord, I am not worthy that thou shouldest come under my roof: but speak the word only, and my servant shall be healed.

et respondens centurio ait Domine non sum dignus ut intres sub tectum meum sed tantum dic uerbo et sanabitur puer

8:9 For I am a man under authority, having soldiers under meus 8:9 nam et ego homo sum sub potestate habens sub me: and I say to this man, Go, and he goeth; and to another, Come, and he cometh; and to my servant, Do this,

and he doeth it. me milites et dico huic uade et uadit et alio ueni et uenit et 8:10 When Jesus heard it, he marvelled, and said to them seruo meo fac hoc et facit 8:10 audiens autem Iesus mira-

that followed, Verily I say unto you, I have not found so great faith, no, not in Israel.

tus est et sequentibus se dixit amen dico uobis non inueni

8:11 And I say unto you, That many shall come from the tantam fidem in Israhel 8:11 dico autem uobis quod multi

east and west, and shall sit down with Abraham, and Isaac, and Jacob, in the kingdom of heaven.

ab oriente et occidente uenient et recumbent cum Abraham

8:12 But the children of the kingdom shall be cast out into et Isaac et Iacob in regno caelorum 8:12 filii autem regni outer darkness: there shall be weeping and gnashing of teeth.

eicientur in tenebras exteriores ibi erit fletus et stridor de-

8:13 And Jesus said unto the centurion, Go thy way; and ntium 8:13 et dixit Iesus centurioni uade et sicut credidisti

as thou hast believed, so be it done unto thee. And his servant was healed in the selfsame hour.

fiat tibi et sanatus est puer in hora illa 8:14 et cum uenisset

8:14 And when Jesus was come into Peter’s house, he saw his wife’s mother laid, and sick of a fever.

Iesus in domum Petri uidit socrum eius iacentem et febrici-

8:15 And he touched her hand, and the fever left her: and tantem 8:15 et tetigit manum eius et dimisit eam febris et

she arose, and ministered unto them.

8:16 When the even was come, they brought unto him surrexit et ministrabat eis 8:16 uespere autem facto obtule-

many that were possessed with devils: and he cast out the spirits with his word, and healed all that were sick:

runt ei multos daemonia habentes et eiciebat spiritus uerbo

8:17 That it might be fulfilled which was spoken by Esaias et omnes male habentes curauit 8:17 ut adimpleretur quod

the prophet, saying, Himself took our infirmities, and bare our sicknesses.

dictum est per Esaiam prophetam dicentem ipse infirmitates

8:18 Now when Jesus saw great multitudes about him, he nostras accepit et aegrotationes portauit 8:18 uidens autem

gave commandment to depart unto the other side.

8:19 And a certain scribe came, and said unto him, Master, Iesus turbas multas circum se iussit ire trans fretum 8:19 et

I will follow thee whithersoever thou goest.

Euangelium Secundum Matthaeum

accedens unus scriba ait illi magister sequar te quocumque ieris 8:20 et dicit ei Iesus uulpes foueas habent et uolucres 8:20 And Jesus saith unto him, The foxes have holes, and

the birds of the air have nests; but the Son of man hath not where to lay his head.

caeli tabernacula Filius autem hominis non habet ubi caput reclinet 8:21 alius autem de discipulis eius ait illi Domine 8:21 And another of his disciples said unto him, Lord, suf-

fer me first to go and bury my father.

permitte me primum ire et sepelire patrem meum 8:22 Ie- 8:22 But Jesus said unto him, Follow me; and let the dead

bury their dead.

sus autem ait illi sequere me et dimitte mortuos sepelire mortuos suos 8:23 et ascendente eo in nauicula secuti sunt eum 8:23 And when he was entered into a ship, his disciples

followed him.

discipuli eius 8:24 et ecce motus magnus factus est in mari 8:24 And, behold, there arose a great tempest in the sea,

insomuch that the ship was covered with the waves: but he was asleep.

ita ut nauicula operiretur fluctibus ipse uero dormiebat 8:25

8:25 And his disciples came to him, and awoke him, saying, Lord, save us: we perish.

et accesserunt et suscitauerunt eum dicentes Domine salua nos perimus 8:26 et dicit eis quid timidi estis modicae fidei 8:26 And he saith unto them, Why are ye fearful, O ye of

little faith? Then he arose, and rebuked the winds and the sea; and there was a great calm.

tunc surgens imperauit uentis et mari et facta est tranquillitas magna 8:27 porro homines mirati sunt dicentes qualis 8:27 But the men marvelled, saying, What manner of man

is this, that even the winds and the sea obey him!

est hic quia et uenti et mare oboediunt ei 8:28 et cum ue- 8:28 And when he was come to the other side into the

country of the Gergesenes, there met him two possessed with devils, coming out of the tombs, exceeding fierce, so

nisset trans fretum in regionem Gerasenorum occurrerunt ei that no man might pass by that way. duo habentes daemonia de monumentis exeuntes saeui nimis ita ut nemo posset transire per uiam illam 8:29 et ecce 8:29 And, behold, they cried out, saying, What have we

to do with thee, Jesus, thou Son of God? art thou come hither to torment us before the time?

clamauerunt dicentes quid nobis et tibi Fili Dei uenisti huc ante tempus torquere nos 8:30 erat autem non longe ab il- 8:30 And there was a good way off from them an herd of

many swine feeding.

lis grex porcorum multorum pascens 8:31 daemones autem 8:31 So the devils besought him, saying, If thou cast us

out, suffer us to go away into the herd of swine.

rogabant eum dicentes si eicis nos mitte nos in gregem porcorum 8:32 et ait illis ite at illi exeuntes abierunt in porcos 8:32 And he said unto them, Go. And when they were

come out, they went into the herd of swine: and, behold, the whole herd of swine ran violently down a steep place

et ecce impetu abiit totus grex per praeceps in mare et mor- into the sea, and perished in the waters. tui sunt in aquis 8:33 pastores autem fugerunt et uenientes 8:33 And they that kept them fled, and went their ways

into the city, and told every thing, and what was befallen to the possessed of the devils.

in ciuitatem nuntiauerunt omnia et de his qui daemonia habuerant 8:34 et ecce tota ciuitas exiit obuiam Iesu et uiso eo 8:34 And, behold, the whole city came out to meet Jesus:

and when they saw him, they besought him that he would depart out of their coasts.

rogabant ut transiret a finibus eorum

9:1 et ascendens in nauiculam transfretauit et uenit in ci- 9:1 And he entered into a ship, and passed over, and came

into his own city.

uitatem suam 9:2 et ecce offerebant ei paralyticum iacen- 9:2 And, behold, they brought to him a man sick of the

palsy, lying on a bed: and Jesus seeing their faith said unto the sick of the palsy; Son, be of good cheer; thy sins

tem in lecto et uidens Iesus fidem illorum dixit paralytico be forgiven thee. confide fili remittuntur tibi peccata tua 9:3 et ecce quidam 9:3 And, behold, certain of the scribes said within them-

selves, This man blasphemeth.

de scribis dixerunt intra se hic blasphemat 9:4 et cum ui- 9:4 And Jesus knowing their thoughts said, Wherefore

think ye evil in your hearts?

disset Iesus cogitationes eorum dixit ut quid cogitatis mala

Biblia Sacra Vulgata

9:5 For whether is easier, to say, Thy sins be forgiven thee; in cordibus uestris 9:5 quid est facilius dicere dimittuntur

or to say, Arise, and walk?

9:6 But that ye may know that the Son of man hath power tibi peccata aut dicere surge et ambula 9:6 ut sciatis autem

on earth to forgive sins, (then saith he to the sick of the palsy,) Arise, take up thy bed, and go unto thine house.

quoniam Filius hominis habet potestatem in terra dimittendi peccata tunc ait paralytico surge tolle lectum tuum et uade in

9:7 And he arose, and departed to his house. domum tuam 9:7 et surrexit et abiit in domum suam 9:8

9:8 But when the multitudes saw it, they marvelled, and glorified God, which had given such power unto men.

uidentes autem turbae timuerunt et glorificauerunt Deum qui

9:9 And as Jesus passed forth from thence, he saw a man, dedit potestatem talem hominibus 9:9 et cum transiret in-

named Matthew, sitting at the receipt of custom: and he saith unto him, Follow me. And he arose, and followed

him. de Iesus uidit hominem sedentem in teloneo Mattheum no-9:10 And it came to pass, as Jesus sat at meat in the house, mine et ait illi sequere me et surgens secutus est eum 9:10

behold, many publicans and sinners came and sat down with him and his disciples.

et factum est discumbente eo in domo ecce multi publicani et peccatores uenientes discumbebant cum Iesu et discipu-

9:11 And when the Pharisees saw it, they said unto his lis eius 9:11 et uidentes Pharisaei dicebant discipulis eius

disciples, Why eateth your Master with publicans and sinners?

quare cum publicanis et peccatoribus manducat magister ue-

9:12 But when Jesus heard that, he said unto them, They ster 9:12 at Iesus audiens ait non est opus ualentibus me-

that be whole need not a physician, but they that are sick.

9:13 But go ye and learn what that meaneth, I will have dico sed male habentibus 9:13 euntes autem discite quid

mercy, and not sacrifice: for I am not come to call the righteous, but sinners to repentance.

est misericordiam uolo et non sacrificium non enim ueni uo-

9:14 Then came to him the disciples of John, saying, Why care iustos sed peccatores 9:14 tunc accesserunt ad eum di-

do we and the Pharisees fast oft, but thy disciples fast not?

scipuli Iohannis dicentes quare nos et Pharisaei ieiunamus

9:15 And Jesus said unto them, Can the children of the frequenter discipuli autem tui non ieiunant 9:15 et ait illis

bridechamber mourn, as long as the bridegroom is with them? but the days will come, when the bridegroom shall

be taken from them, and then shall they fast. Iesus numquid possunt filii sponsi lugere quamdiu cum illis

est sponsus uenient autem dies cum auferetur ab eis sponsus

9:16 No man putteth a piece of new cloth unto an old gar- et tunc ieiunabunt 9:16 nemo autem inmittit commissuram

ment, for that which is put in to fill it up taketh from the garment, and the rent is made worse.

panni rudis in uestimentum uetus tollit enim plenitudinem

9:17 Neither do men put new wine into old bottles: else eius a uestimento et peior scissura fit 9:17 neque mittunt

the bottles break, and the wine runneth out, and the bottles perish: but they put new wine into new bottles, and both

are preserved. uinum nouum in utres ueteres alioquin rumpuntur utres et

uinum effunditur et utres pereunt sed uinum nouum in utres

9:18 While he spake these things unto them, behold, there nouos mittunt et ambo conseruantur 9:18 haec illo loquente

came a certain ruler, and worshipped him, saying, My daughter is even now dead: but come and lay thy hand

upon her, and she shall live. ad eos ecce princeps unus accessit et adorabat eum dicens

filia mea modo defuncta est sed ueni inpone manum super

9:19 And Jesus arose, and followed him, and so did his eam et uiuet 9:19 et surgens Iesus sequebatur eum et disci-

disciples.

9:20 And, behold, a woman, which was diseased with an puli eius 9:20 et ecce mulier quae sanguinis fluxum patie-

issue of blood twelve years, came behind him, and touched the hem of his garment:

batur duodecim annis accessit retro et tetigit fimbriam uesti-

9:21 For she said within herself, If I may but touch his menti eius 9:21 dicebat enim intra se si tetigero tantum ue-

garment, I shall be whole.

Euangelium Secundum Matthaeum

stimentum eius salua ero 9:22 at Iesus conuersus et uidens 9:22 But Jesus turned him about, and when he saw her, he

said, Daughter, be of good comfort; thy faith hath made thee whole. And the woman was made whole from that

eam dixit confide filia fides tua te saluam fecit et salua facta hour. est mulier ex illa hora 9:23 et cum uenisset Iesus in domum 9:23 And when Jesus came into the ruler’s house, and saw

the minstrels and the people making a noise,

principis et uidisset tibicines et turbam tumultuantem 9:24 9:24 He said unto them, Give place: for the maid is not

dead, but sleepeth. And they laughed him to scorn.

dicebat recedite non est enim mortua puella sed dormit et deridebant eum 9:25 et cum eiecta esset turba intrauit et tenuit 9:25 But when the people were put forth, he went in, and

took her by the hand, and the maid arose.

manum eius et surrexit puella 9:26 et exiit fama haec in uni- 9:26 And the fame hereof went abroad into all that land. uersam terram illam 9:27 et transeunte inde Iesu secuti sunt 9:27 And when Jesus departed thence, two blind men fol-

lowed him, crying, and saying, Thou son of David, have mercy on us.

eum duo caeci clamantes et dicentes miserere nostri Fili Dauid 9:28 cum autem uenisset domum accesserunt ad eum 9:28 And when he was come into the house, the blind men

came to him: and Jesus saith unto them, Believe ye that I am able to do this? They said unto him, Yea, Lord.

caeci et dicit eis Iesus creditis quia possum hoc facere uobis dicunt ei utique Domine 9:29 tunc tetigit oculos eorum di- 9:29 Then touched he their eyes, saying, According to

your faith be it unto you.

cens secundum fidem uestram fiat uobis 9:30 et aperti sunt 9:30 And their eyes were opened; and Jesus straitly

charged them, saying, See that no man know it.

oculi illorum et comminatus est illis Iesus dicens uidete ne quis sciat 9:31 illi autem exeuntes diffamauerunt eum in 9:31 But they, when they were departed, spread abroad his

fame in all that country.

tota terra illa 9:32 egressis autem illis ecce obtulerunt ei 9:32 As they went out, behold, they brought to him a dumb

man possessed with a devil.

hominem mutum daemonium habentem 9:33 et eiecto dae- 9:33 And when the devil was cast out, the dumb spake:

and the multitudes marvelled, saying, It was never so seen in Israel.

mone locutus est mutus et miratae sunt turbae dicentes numquam paruit sic in Israhel 9:34 Pharisaei autem dicebant in 9:34 But the Pharisees said, He casteth out devils through

the prince of the devils.

principe daemoniorum eicit daemones 9:35 et circumibat 9:35 And Jesus went about all the cities and villages,

teaching in their synagogues, and preaching the gospel of the kingdom, and healing every sickness and every disease

Iesus ciuitates omnes et castella docens in synagogis eorum among the people. et praedicans euangelium regni et curans omnem languorem et omnem infirmitatem 9:36 uidens autem turbas misertus 9:36 But when he saw the multitudes, he was moved with

compassion on them, because they fainted, and were scattered abroad, as sheep having no shepherd.

est eis quia erant uexati et iacentes sicut oues non habentes pastorem 9:37 tunc dicit discipulis suis messis quidem 9:37 Then saith he unto his disciples, The harvest truly is

plenteous, but the labourers are few;

multa operarii autem pauci 9:38 rogate ergo dominum mes- 9:38 Pray ye therefore the Lord of the harvest, that he will

send forth labourers into his harvest.

sis ut eiciat operarios in messem suam

10:1 et conuocatis duodecim discipulis suis dedit illis po- 10:1 And when he had called unto him his twelve disci-

ples, he gave them power against unclean spirits, to cast them out, and to heal all manner of sickness and all man-

testatem spirituum inmundorum ut eicerent eos et curarent ner of disease.

omnem languorem et omnem infirmitatem 10:2 duodecim 10:2 Now the names of the twelve apostles are these; The first, Simon, who is called Peter, and Andrew his brother; James the son of Zebedee, and John his brother;

autem apostolorum nomina sunt haec primus Simon qui dicitur Petrus et Andreas frater eius 10:3 Iacobus Zebedaei 10:3 Philip, and Bartholomew; Thomas, and Matthew

the publican; James the son of Alphaeus, and Lebbaeus, whose surname was Thaddaeus;

et Iohannes frater eius Philippus et Bartholomeus Thomas

Biblia Sacra Vulgata

10:4 Simon the Canaanite, and Judas Iscariot, who also et Mattheus publicanus et Iacobus Alphei et Thaddeus 10:4

betrayed him.

10:5 These twelve Jesus sent forth, and commanded them, Simon Cananeus et Iudas Scariotes qui et tradidit eum 10:5

saying, Go not into the way of the Gentiles, and into any city of the Samaritans enter ye not:

hos duodecim misit Iesus praecipiens eis et dicens in uiam gentium ne abieritis et in ciuitates Samaritanorum ne intra-

10:6 But go rather to the lost sheep of the house of Israel. ueritis 10:6 sed potius ite ad oues quae perierunt domus Is-10:7 And as ye go, preach, saying, The kingdom of heaven rahel 10:7 euntes autem praedicate dicentes quia adpropin-

is at hand. 10:8 Heal the sick, cleanse the lepers, raise the dead, cast quauit regnum caelorum 10:8 infirmos curate mortuos suout devils: freely ye have received, freely give.

scitate leprosos mundate daemones eicite gratis accepistis

10:9 Provide neither gold, nor silver, nor brass in your gratis date 10:9 nolite possidere aurum neque argentum ne-

purses, 10:10 Nor scrip for your journey, neither two coats, neither que pecuniam in zonis uestris 10:10 non peram in uia neque shoes, nor yet staves: for the workman is worthy of his meat.

duas tunicas neque calciamenta neque uirgam dignus enim

10:11 And into whatsoever city or town ye shall enter, en- est operarius cibo suo 10:11 in quamcumque ciuitatem aut quire who in it is worthy; and there abide till ye go thence.

castellum intraueritis interrogate quis in ea dignus sit et ibi

10:12 And when ye come into an house, salute it. manete donec exeatis 10:12 intrantes autem in domum sa-10:13 And if the house be worthy, let your peace come lutate eam 10:13 et siquidem fuerit domus digna ueniat pax

upon it: but if it be not worthy, let your peace return to you.

uestra super eam si autem non fuerit digna pax uestra ad uos

10:14 And whosoever shall not receive you, nor hear your reuertatur 10:14 et quicumque non receperit uos neque au-

words, when ye depart out of that house or city, shake off the dust of your feet.

dierit sermones uestros exeuntes foras de domo uel de ciui-

10:15 Verily I say unto you, It shall be more tolerable for tate excutite puluerem de pedibus uestris 10:15 amen dico

the land of Sodom and Gomorrha in the day of judgment, than for that city.

uobis tolerabilius erit terrae Sodomorum et Gomorraeorum

10:16 Behold, I send you forth as sheep in the midst of in die iudicii quam illi ciuitati 10:16 ecce ego mitto uos sic-

wolves: be ye therefore wise as serpents, and harmless as doves.

ut oues in medio luporum estote ergo prudentes sicut ser-

10:17 But beware of men: for they will deliver you up pentes et simplices sicut columbae 10:17 cauete autem ab

to the councils, and they will scourge you in their synagogues;

hominibus tradent enim uos in conciliis et in synagogis suis

10:18 And ye shall be brought before governors and kings flagellabunt uos 10:18 et ad praesides et ad reges ducemini

for my sake, for a testimony against them and the Gentiles.

10:19 But when they deliver you up, take no thought how propter me in testimonium illis et gentibus 10:19 cum au-

or what ye shall speak: for it shall be given you in that same hour what ye shall speak.

tem tradent uos nolite cogitare quomodo aut quid loquamini

10:20 For it is not ye that speak, but the Spirit of your dabitur enim uobis in illa hora quid loquamini 10:20 non

Father which speaketh in you.

enim uos estis qui loquimini sed Spiritus Patris uestri qui lo-

10:21 And the brother shall deliver up the brother to death, quitur in uobis 10:21 tradet autem frater fratrem in mortem and the father the child: and the children shall rise up against their parents, and cause them to be put to death.

et pater filium et insurgent filii in parentes et morte eos ad-

10:22 And ye shall be hated of all men for my name’s sake: ficient 10:22 et eritis odio omnibus propter nomen meum but he that endureth to the end shall be saved.

10:23 But when they persecute you in this city, flee ye into qui autem perseuerauerit in finem hic saluus erit 10:23 cum

another: for verily I say unto you, Ye shall not have gone over the cities of Israel, till the Son of man be come.

Euangelium Secundum Matthaeum

autem persequentur uos in ciuitate ista fugite in aliam amen enim dico uobis non consummabitis ciuitates Israhel donec ueniat Filius hominis 10:24 non est discipulus super magi- 10:24 The disciple is not above his master, nor the servant

above his lord.

strum nec seruus super dominum suum 10:25 sufficit disci- 10:25 It is enough for the disciple that he be as his master, and the servant as his lord. If they have called the master of the house Beelzebub, how much more shall they call

pulo ut sit sicut magister eius et seruus sicut dominus eius si them of his household? patrem familias Beelzebub uocauerunt quanto magis domesticos eius 10:26 ne ergo timueritis eos nihil enim opertum 10:26 Fear them not therefore: for there is nothing cov-

ered, that shall not be revealed; and hid, that shall not be known.

quod non reuelabitur et occultum quod non scietur 10:27

10:27 What I tell you in darkness, that speak ye in light: and what ye hear in the ear, that preach ye upon the housetops.

quod dico uobis in tenebris dicite in lumine et quod in aure auditis praedicate super tecta 10:28 et nolite timere eos qui 10:28 And fear not them which kill the body, but are not

able to kill the soul: but rather fear him which is able to destroy both soul and body in hell.

occidunt corpus animam autem non possunt occidere sed potius eum timete qui potest et animam et corpus perdere in gehennam 10:29 nonne duo passeres asse ueneunt et unus ex 10:29 Are not two sparrows sold for a farthing? and one

of them shall not fall on the ground without your Father.

illis non cadet super terram sine Patre uestro 10:30 uestri 10:30 But the very hairs of your head are all numbered. autem et capilli capitis omnes numerati sunt 10:31 nolite 10:31 Fear ye not therefore, ye are of more value than

many sparrows.

ergo timere multis passeribus meliores estis uos 10:32 om- 10:32 Whosoever therefore shall confess me before men,

him will I confess also before my Father which is in heaven.

nis ergo qui confitebitur me coram hominibus confitebor et ego eum coram Patre meo qui est in caelis 10:33 qui autem 10:33 But whosoever shall deny me before men, him will

I also deny before my Father which is in heaven.

negauerit me coram hominibus negabo et ego eum coram Patre meo qui est in caelis 10:34 nolite arbitrari quia ue- 10:34 Think not that I am come to send peace on earth: I

came not to send peace, but a sword.

nerim mittere pacem in terram non ueni pacem mittere sed gladium 10:35 ueni enim separare hominem aduersus pat- 10:35 For I am come to set a man at variance against his

father, and the daughter against her mother, and the daughter in law against her mother in law.

rem suum et filiam aduersus matrem suam et nurum aduersus socrum suam 10:36 et inimici hominis domestici eius 10:36 And a man’s foes shall be they of his own house-

hold.

10:37 qui amat patrem aut matrem plus quam me non est me 10:37 He that loveth father or mother more than me is not worthy of me: and he that loveth son or daughter more than me is not worthy of me.

dignus et qui amat filium aut filiam super me non est me di-

gnus 10:38 et qui non accipit crucem suam et sequitur me 10:38 And he that taketh not his cross, and followeth after

me, is not worthy of me.

non est me dignus 10:39 qui inuenit animam suam perdet il- 10:39 He that findeth his life shall lose it: and he that los-

eth his life for my sake shall find it.

lam et qui perdiderit animam suam propter me inueniet eam 10:40 qui recipit uos me recipit et qui me recipit recipit eum 10:40 He that receiveth you receiveth me, and he that re-

ceiveth me receiveth him that sent me.

qui me misit 10:41 qui recipit prophetam in nomine pro- 10:41 He that receiveth a prophet in the name of a prophet shall receive a prophet’s reward; and he that receiveth a righteous man in the name of a righteous man shall receive

phetae mercedem prophetae accipiet et qui recipit iustum in a righteous man’s reward. nomine iusti mercedem iusti accipiet 10:42 et quicumque 10:42 And whosoever shall give to drink unto one of these

little ones a cup of cold water only in the name of a disciple, verily I say unto you, he shall in no wise lose his reward.

Biblia Sacra Vulgata

potum dederit uni ex minimis istis calicem aquae frigidae tantum in nomine discipuli amen dico uobis non perdet mercedem suam

11:1 And it came to pass, when Jesus had made an end of 11:1 et factum est cum consummasset Iesus praecipiens

commanding his twelve disciples, he departed thence to teach and to preach in their cities.

duodecim discipulis suis transiit inde ut doceret et praedi-

11:2 Now when John had heard in the prison the works of caret in ciuitatibus eorum 11:2 Iohannes autem cum audis-

Christ, he sent two of his disciples,

set in uinculis opera Christi mittens duos de discipulis suis

11:3 And said unto him, Art thou he that should come, or 11:3 ait illi tu es qui uenturus es an alium expectamus 11:4

do we look for another? 11:4 Jesus answered and said unto them, Go and shew John again those things which ye do hear and see:

et respondens Iesus ait illis euntes renuntiate Iohanni quae

11:5 The blind receive their sight, and the lame walk, the auditis et uidetis 11:5 caeci uident claudi ambulant leprosi

lepers are cleansed, and the deaf hear, the dead are raised up, and the poor have the gospel preached to them.

mundantur surdi audiunt mortui resurgunt pauperes euange-

11:6 And blessed is he, whosoever shall not be offended lizantur 11:6 et beatus est qui non fuerit scandalizatus in

in me.

11:7 And as they departed, Jesus began to say unto the me 11:7 illis autem abeuntibus coepit Iesus dicere ad tur-

multitudes concerning John, What went ye out into the wilderness to see? A reed shaken with the wind?

bas de Iohanne quid existis in desertum uidere harundinem

11:8 But what went ye out for to see? A man clothed in uento agitatam 11:8 sed quid existis uidere hominem molli-

soft raiment? behold, they that wear soft clothing are in kings’ houses.

bus uestitum ecce qui mollibus uestiuntur in domibus regum

11:9 But what went ye out for to see? A prophet? yea, I sunt 11:9 sed quid existis uidere prophetam etiam dico uo-

say unto you, and more than a prophet. 11:10 For this is he, of whom it is written, Behold, I send bis et plus quam prophetam 11:10 hic enim est de quo scmy messenger before thy face, which shall prepare thy way before thee.

riptum est ecce ego mitto angelum meum ante faciem tuam

11:11 Verily I say unto you, Among them that are born of qui praeparabit uiam tuam ante te 11:11 amen dico uobis

women there hath not risen a greater than John the Baptist: notwithstanding he that is least in the kingdom of heaven

is greater than he. non surrexit inter natos mulierum maior Iohanne Baptista 11:12 And from the days of John the Baptist until now the qui autem minor est in regno caelorum maior est illo 11:12

kingdom of heaven suffereth violence, and the violent take it by force.

a diebus autem Iohannis Baptistae usque nunc regnum caelo-

11:13 For all the prophets and the law prophesied until rum uim patitur et uiolenti rapiunt illud 11:13 omnes enim

John. 11:14 And if ye will receive it, this is Elias, which was for prophetae et lex usque ad Iohannem prophetauerunt 11:14 to come.

11:15 He that hath ears to hear, let him hear. et si uultis recipere ipse est Helias qui uenturus est 11:15

11:16 But whereunto shall I liken this generation? It is qui habet aures audiendi audiat 11:16 cui autem similem like unto children sitting in the markets, and calling unto their fellows,

aestimabo generationem istam similis est pueris sedentibus

11:17 And saying, We have piped unto you, and ye have in foro qui clamantes coaequalibus 11:17 dicunt cecinimus

not danced; we have mourned unto you, and ye have not lamented.

uobis et non saltastis lamentauimus et non planxistis 11:18

11:18 For John came neither eating nor drinking, and they say, He hath a devil.

uenit enim Iohannes neque manducans neque bibens et di-

11:19 The Son of man came eating and drinking, and they cunt daemonium habet 11:19 uenit Filius hominis mandu-

say, Behold a man gluttonous, and a winebibber, a friend of publicans and sinners. But wisdom is justified of her

children. cans et bibens et dicunt ecce homo uorax et potator uini pub-

Euangelium Secundum Matthaeum

licanorum et peccatorum amicus et iustificata est sapientia a filiis suis 11:20 tunc coepit exprobrare ciuitatibus in quibus 11:20 Then began he to upbraid the cities wherein most of

his mighty works were done, because they repented not:

factae sunt plurimae uirtutes eius quia non egissent paenitentiam 11:21 uae tibi Corazain uae tibi Bethsaida quia si 11:21 Woe unto thee, Chorazin! woe unto thee, Bethsaida!

for if the mighty works, which were done in you, had been done in Tyre and Sidon, they would have repented long

in Tyro et Sidone factae essent uirtutes quae factae sunt in ago in sackcloth and ashes. uobis olim in cilicio et cinere paenitentiam egissent 11:22 11:22 But I say unto you, It shall be more tolerable for

Tyre and Sidon at the day of judgment, than for you.

uerumtamen dico uobis Tyro et Sidoni remissius erit in die iudicii quam uobis 11:23 et tu Capharnaum numquid usque 11:23 And thou, Capernaum, which art exalted unto

heaven, shalt be brought down to hell: for if the mighty works, which have been done in thee, had been done in

in caelum exaltaberis usque in infernum descendes quia si in Sodom, it would have remained until this day. Sodomis factae fuissent uirtutes quae factae sunt in te forte mansissent usque in hunc diem 11:24 uerumtamen dico uo- 11:24 But I say unto you, That it shall be more tolerable

for the land of Sodom in the day of judgment, than for thee.

bis quia terrae Sodomorum remissius erit in die iudicii quam tibi 11:25 in illo tempore respondens Iesus dixit confiteor 11:25 At that time Jesus answered and said, I thank thee,

O Father, Lord of heaven and earth, because thou hast hid these things from the wise and prudent, and hast revealed

tibi Pater Domine caeli et terrae quia abscondisti haec a sa- them unto babes. pientibus et prudentibus et reuelasti ea paruulis 11:26 ita 11:26 Even so, Father: for so it seemed good in thy sight. Pater quoniam sic fuit placitum ante te 11:27 omnia mihi 11:27 All things are delivered unto me of my Father: and

no man knoweth the Son, but the Father; neither knoweth any man the Father, save the Son, and he to whomsoever

tradita sunt a Patre meo et nemo nouit Filium nisi Pater ne- the Son will reveal him. que Patrem quis nouit nisi Filius et cui uoluerit Filius reuelare 11:28 uenite ad me omnes qui laboratis et onerati estis 11:28 Come unto me, all ye that labour and are heavy

laden, and I will give you rest.

et ego reficiam uos 11:29 tollite iugum meum super uos et 11:29 Take my yoke upon you, and learn of me; for I am

meek and lowly in heart: and ye shall find rest unto your souls.

discite a me quia mitis sum et humilis corde et inuenietis requiem animabus uestris 11:30 iugum enim meum suaue est 11:30 For my yoke is easy, and my burden is light. et onus meum leue est

12:1 in illo tempore abiit Iesus sabbato per sata discipuli 12:1 At that time Jesus went on the sabbath day through

the corn; and his disciples were an hungred, and began to pluck the ears of corn and to eat.

autem eius esurientes coeperunt uellere spicas et manducare 12:2 Pharisaei autem uidentes dixerunt ei ecce discipuli tui 12:2 But when the Pharisees saw it, they said unto him,

Behold, thy disciples do that which is not lawful to do upon the sabbath day.

faciunt quod non licet eis facere sabbatis 12:3 at ille dixit

12:3 But he said unto them, Have ye not read what David did, when he was an hungred, and they that were with him;

eis non legistis quid fecerit Dauid quando esuriit et qui cum eo erant 12:4 quomodo intrauit in domum Dei et panes pro- 12:4 How he entered into the house of God, and did eat the

shewbread, which was not lawful for him to eat, neither for them which were with him, but only for the priests?

positionis comedit quos non licebat ei edere neque his qui cum eo erant nisi solis sacerdotibus 12:5 aut non legistis in 12:5 Or have ye not read in the law, how that on the sab-

bath days the priests in the temple profane the sabbath, and are blameless?

lege quia sabbatis sacerdotes in templo sabbatum uiolant et sine crimine sunt 12:6 dico autem uobis quia templo maior 12:6 But I say unto you, That in this place is one greater

than the temple.

Biblia Sacra Vulgata

12:7 But if ye had known what this meaneth, I will have est hic 12:7 si autem sciretis quid est misericordiam uolo et

mercy, and not sacrifice, ye would not have condemned the guiltless.

non sacrificium numquam condemnassetis innocentes 12:8

12:8 For the Son of man is Lord even of the sabbath day.

12:9 And when he was departed thence, he went into their dominus est enim Filius hominis etiam sabbati 12:9 et cum

synagogue:

12:10 And, behold, there was a man which had his hand inde transisset uenit in synagogam eorum 12:10 et ecce

withered. And they asked him, saying, Is it lawful to heal on the sabbath days? that they might accuse him.

homo manum habens aridam et interrogabant eum dicentes

12:11 And he said unto them, What man shall there be si licet sabbatis curare ut accusarent eum 12:11 ipse autem

among you, that shall have one sheep, and if it fall into a pit on the sabbath day, will he not lay hold on it, and lift it

out? dixit illis quis erit ex uobis homo qui habeat ouem unam et

si ceciderit haec sabbatis in foueam nonne tenebit et leuabit

12:12 How much then is a man better than a sheep? eam 12:12 quanto magis melior est homo oue itaque licet

Wherefore it is lawful to do well on the sabbath days.

12:13 Then saith he to the man, Stretch forth thine hand. sabbatis benefacere 12:13 tunc ait homini extende manum

And he stretched it forth; and it was restored whole, like as the other.

tuam et extendit et restituta est sanitati sicut altera 12:14

12:14 Then the Pharisees went out, and held a council against him, how they might destroy him.

exeuntes autem Pharisaei consilium faciebant aduersus eum

12:15 But when Jesus knew it, he withdrew himself from quomodo eum perderent 12:15 Iesus autem sciens recessit

thence: and great multitudes followed him, and he healed them all;

inde et secuti sunt eum multi et curauit eos omnes 12:16

12:16 And charged them that they should not make him known:

12:17 That it might be fulfilled which was spoken by Esa- et praecepit eis ne manifestum eum facerent 12:17 ut ad-

ias the prophet, saying,

impleretur quod dictum est per Esaiam prophetam dicentem

12:18 Behold my servant, whom I have chosen; my 12:18 ecce puer meus quem elegi dilectus meus in quo bene

beloved, in whom my soul is well pleased: I will put my spirit upon him, and he shall shew judgment to the Gen-

tiles. placuit animae meae ponam spiritum meum super eum et 12:19 He shall not strive, nor cry; neither shall any man iudicium gentibus nuntiabit 12:19 non contendet neque cla-

hear his voice in the streets.

12:20 A bruised reed shall he not break, and smoking flax mabit neque audiet aliquis in plateis uocem eius 12:20 ha-

shall he not quench, till he send forth judgment unto victory.

rundinem quassatam non confringet et linum fumigans non

12:21 And in his name shall the Gentiles trust. extinguet donec eiciat ad uictoriam iudicium 12:21 et in no-12:22 Then was brought unto him one possessed with a mine eius gentes sperabunt 12:22 tunc oblatus est ei daemo-

devil, blind, and dumb: and he healed him, insomuch that the blind and dumb both spake and saw.

nium habens caecus et mutus et curauit eum ita ut loqueretur

12:23 And all the people were amazed, and said, Is not et uideret 12:23 et stupebant omnes turbae et dicebant num-

this the son of David? 12:24 But when the Pharisees heard it, they said, This fel- quid hic est Filius Dauid 12:24 Pharisaei autem audientes low doth not cast out devils, but by Beelzebub the prince of the devils.

dixerunt hic non eicit daemones nisi in Beelzebub principe

12:25 And Jesus knew their thoughts, and said unto them, daemoniorum 12:25 Iesus autem sciens cogitationes eorum

Every kingdom divided against itself is brought to desolation; and every city or house divided against itself shall

not stand: dixit eis omne regnum diuisum contra se desolatur et om-12:26 And if Satan cast out Satan, he is divided against nis ciuitas uel domus diuisa contra se non stabit 12:26 et si

himself; how shall then his kingdom stand?

Satanas Satanan eicit aduersus se diuisus est quomodo ergo

12:27 And if I by Beelzebub cast out devils, by whom do stabit regnum eius 12:27 et si ego in Beelzebub eicio dae-

your children cast them out? therefore they shall be your judges.

mones filii uestri in quo eiciunt ideo ipsi iudices erunt uestri

Euangelium Secundum Matthaeum

12:28 si autem ego in Spiritu Dei eicio daemones igitur perue- 12:28 But if I cast out devils by the Spirit of God, then the

kingdom of God is come unto you.

nit in uos regnum Dei 12:29 aut quomodo potest quisquam 12:29 Or else how can one enter into a strong man’s house, and spoil his goods, except he first bind the strong man? and then he will spoil his house.

intrare in domum fortis et uasa eius diripere nisi prius alligauerit fortem et tunc domum illius diripiat 12:30 qui non est 12:30 He that is not with me is against me; and he that

gathereth not with me scattereth abroad.

mecum contra me est et qui non congregat mecum spargit 12:31 ideo dico uobis omne peccatum et blasphemia remit- 12:31 Wherefore I say unto you, All manner of sin and

blasphemy shall be forgiven unto men: but the blasphemy against the Holy Ghost shall not be forgiven unto men.

tetur hominibus Spiritus autem blasphemia non remittetur 12:32 et quicumque dixerit uerbum contra Filium hominis re- 12:32 And whosoever speaketh a word against the Son

of man, it shall be forgiven him: but whosoever speaketh against the Holy Ghost, it shall not be forgiven him, nei-

mittetur ei qui autem dixerit contra Spiritum Sanctum non ther in this world, neither in the world to come. remittetur ei neque in hoc saeculo neque in futuro 12:33 aut 12:33 Either make the tree good, and his fruit good; or else

make the tree corrupt, and his fruit corrupt: for the tree is known by his fruit.

facite arborem bonam et fructum eius bonum aut facite arborem malam et fructum eius malum siquidem ex fructu arbor agnoscitur 12:34 progenies uiperarum quomodo potes- 12:34 O generation of vipers, how can ye, being evil,

speak good things? for out of the abundance of the heart the mouth speaketh.

tis bona loqui cum sitis mali ex abundantia enim cordis os loquitur 12:35 bonus homo de bono thesauro profert bona et 12:35 A good man out of the good treasure of the heart

bringeth forth good things: and an evil man out of the evil treasure bringeth forth evil things.

malus homo de malo thesauro profert mala 12:36 dico au-

12:36 But I say unto you, That every idle word that men shall speak, they shall give account thereof in the day of judgment.

tem uobis quoniam omne uerbum otiosum quod locuti fuerint homines reddent rationem de eo in die iudicii 12:37 ex 12:37 For by thy words thou shalt be justified, and by thy

words thou shalt be condemned.

uerbis enim tuis iustificaberis et ex uerbis tuis condemnaberis 12:38 tunc responderunt ei quidam de scribis et Phari- 12:38 Then certain of the scribes and of the Pharisees an-

swered, saying, Master, we would see a sign from thee.

saeis dicentes magister uolumus a te signum uidere 12:39 12:39 But he answered and said unto them, An evil and

adulterous generation seeketh after a sign; and there shall no sign be given to it, but the sign of the prophet Jonas:

qui respondens ait illis generatio mala et adultera signum quaerit et signum non dabitur ei nisi signum Ionae prophetae 12:40 sicut enim fuit Ionas in uentre ceti tribus diebus 12:40 For as Jonas was three days and three nights in the

whale’s belly; so shall the Son of man be three days and three nights in the heart of the earth.

et tribus noctibus sic erit Filius hominis in corde terrae tribus diebus et tribus noctibus 12:41 uiri nineuitae surgent in 12:41 The men of Nineveh shall rise in judgment with this

generation, and shall condemn it: because they repented at the preaching of Jonas; and, behold, a greater than Jonas

iudicio cum generatione ista et condemnabunt eam quia pae- is here. nitentiam egerunt in praedicatione Ionae et ecce plus quam Iona hic 12:42 regina austri surget in iudicio cum genera- 12:42 The queen of the south shall rise up in the judgment

with this generation, and shall condemn it: for she came from the uttermost parts of the earth to hear the wisdom of

tione ista et condemnabit eam quia uenit a finibus terrae au- Solomon; and, behold, a greater than Solomon is here. dire sapientiam Salomonis et ecce plus quam Salomon hic 12:43 cum autem inmundus spiritus exierit ab homine am- 12:43 When the unclean spirit is gone out of a man, he

walketh through dry places, seeking rest, and findeth none.

bulat per loca arida quaerens requiem et non inuenit 12:44 12:44 Then he saith, I will return into my house from

whence I came out; and when he is come, he findeth it empty, swept, and garnished.

Biblia Sacra Vulgata

tunc dicit reuertar in domum meam unde exiui et ueniens

12:45 Then goeth he, and taketh with himself seven other inuenit uacantem scopis mundatam et ornatam 12:45 tunc spirits more wicked than himself, and they enter in and dwell there: and the last state of that man is worse than the

first. Even so shall it be also unto this wicked generation. uadit et adsumit septem alios spiritus secum nequiores se et

intrantes habitant ibi et fiunt nouissima hominis illius peiora

12:46 While he yet talked to the people, behold, his prioribus sic erit et generationi huic pessimae 12:46 adhuc

mother and his brethren stood without, desiring to speak with him.

eo loquente ad turbas ecce mater eius et fratres stabant foris

12:47 Then one said unto him, Behold, thy mother and thy quaerentes loqui ei 12:47 dixit autem ei quidam ecce mater

brethren stand without, desiring to speak with thee.

12:48 But he answered and said unto him that told him, tua et fratres tui foris stant quaerentes te 12:48 at ipse res-

Who is my mother? and who are my brethren?

pondens dicenti sibi ait quae est mater mea et qui sunt fratres

12:49 And he stretched forth his hand toward his disciples, mei 12:49 et extendens manum in discipulos suos dixit ecce and said, Behold my mother and my brethren!

12:50 For whosoever shall do the will of my Father which mater mea et fratres mei 12:50 quicumque enim fecerit uo-

is in heaven, the same is my brother, and sister, and mother.

luntatem Patris mei qui in caelis est ipse meus et frater et soror et mater est

13:1 The same day went Jesus out of the house, and sat by 13:1 in illo die exiens Iesus de domo sedebat secus mare

the sea side.

13:2 And great multitudes were gathered together unto 13:2 et congregatae sunt ad eum turbae multae ita ut in naui-

him, so that he went into a ship, and sat; and the whole multitude stood on the shore.

culam ascendens sederet et omnis turba stabat in litore 13:3

13:3 And he spake many things unto them in parables, saying, Behold, a sower went forth to sow;

et locutus est eis multa in parabolis dicens ecce exiit qui se-

13:4 And when he sowed, some seeds fell by the way side, minat seminare 13:4 et dum seminat quaedam ceciderunt

and the fowls came and devoured them up:

13:5 Some fell upon stony places, where they had not secus uiam et uenerunt uolucres et comederunt ea 13:5 alia

much earth: and forthwith they sprung up, because they had no deepness of earth:

autem ceciderunt in petrosa ubi non habebat terram multam et continuo exorta sunt quia non habebant altitudinem

13:6 And when the sun was up, they were scorched; and terrae 13:6 sole autem orto aestuauerunt et quia non habe-

because they had no root, they withered away.

13:7 And some fell among thorns; and the thorns sprung bant radicem aruerunt 13:7 alia autem ceciderunt in spinas

up, and choked them:

13:8 But other fell into good ground, and brought forth et creuerunt spinae et suffocauerunt ea 13:8 alia uero ce-

fruit, some an hundredfold, some sixtyfold, some thirtyfold.

ciderunt in terram bonam et dabant fructum aliud centesi-

13:9 Who hath ears to hear, let him hear. mum aliud sexagesimum aliud tricesimum 13:9 qui habet 13:10 And the disciples came, and said unto him, Why aures audiendi audiat 13:10 et accedentes discipuli dixerunt

speakest thou unto them in parables?

13:11 He answered and said unto them, Because it is given ei quare in parabolis loqueris eis 13:11 qui respondens ait

unto you to know the mysteries of the kingdom of heaven, but to them it is not given.

illis quia uobis datum est nosse mysteria regni caelorum il-

13:12 For whosoever hath, to him shall be given, and he lis autem non est datum 13:12 qui enim habet dabitur ei et

shall have more abundance: but whosoever hath not, from him shall be taken away even that he hath.

abundabit qui autem non habet et quod habet auferetur ab eo

13:13 Therefore speak I to them in parables: because they 13:13 ideo in parabolis loquor eis quia uidentes non uident et

seeing see not; and hearing they hear not, neither do they understand.

audientes non audiunt neque intellegunt 13:14 et adimple-

13:14 And in them is fulfilled the prophecy of Esaias, which saith, By hearing ye shall hear, and shall not understand; and seeing ye shall see, and shall not perceive:

Euangelium Secundum Matthaeum

tur eis prophetia Esaiae dicens auditu audietis et non intellegetis et uidentes uidebitis et non uidebitis 13:15 incrassa- 13:15 For this people’s heart is waxed gross, and their ears

are dull of hearing, and their eyes they have closed; lest at any time they should see with their eyes and hear with their ears, and should understand with their heart, and should be converted, and I should heal them.

tum est enim cor populi huius et auribus grauiter audierunt et oculos suos cluserunt nequando oculis uideant et auribus audiant et corde intellegant et conuertantur et sanem eos 13:16 13:16 But blessed are your eyes, for they see: and your

ears, for they hear.

uestri autem beati oculi quia uident et aures uestrae quia audiunt 13:17 amen quippe dico uobis quia multi prophetae et 13:17 For verily I say unto you, That many prophets and

righteous men have desired to see those things which ye see, and have not seen them; and to hear those things

iusti cupierunt uidere quae uidetis et non uiderunt et audire which ye hear, and have not heard them. quae auditis et non audierunt 13:18 uos ergo audite parabo- 13:18 Hear ye therefore the parable of the sower. lam seminantis 13:19 omnis qui audit uerbum regni et non 13:19 When any one heareth the word of the kingdom,

and understandeth it not, then cometh the wicked one, and catcheth away that which was sown in his heart. This is he

intellegit uenit malus et rapit quod seminatum est in corde which received seed by the way side.

eius hic est qui secus uiam seminatus est 13:20 qui autem 13:20 But he that received the seed into stony places, the same is he that heareth the word, and anon with joy receiveth it;

supra petrosa seminatus est hic est qui uerbum audit et continuo cum gaudio accipit illud 13:21 non habet autem in se 13:21 Yet hath he not root in himself, but dureth for a

while: for when tribulation or persecution ariseth because of the word, by and by he is offended.

radicem sed est temporalis facta autem tribulatione et persecutione propter uerbum continuo scandalizatur 13:22 qui 13:22 He also that received seed among the thorns is he

that heareth the word; and the care of this world, and the deceitfulness of riches, choke the word, and he becometh

autem est seminatus in spinis hic est qui uerbum audit et sol- unfruitful. licitudo saeculi istius et fallacia diuitiarum suffocat uerbum et sine fructu efficitur 13:23 qui uero in terra bona semina- 13:23 But he that received seed into the good ground is he

that heareth the word, and understandeth it; which also beareth fruit, and bringeth forth, some an hundredfold,

tus est hic est qui audit uerbum et intellegit et fructum adfert some sixty, some thirty. et facit aliud quidem centum aliud autem sexaginta porro aliud triginta 13:24 aliam parabolam proposuit illis dicens 13:24 Another parable put he forth unto them, saying, The

kingdom of heaven is likened unto a man which sowed good seed in his field:

simile factum est regnum caelorum homini qui seminauit bonum semen in agro suo 13:25 cum autem dormirent homi- 13:25 But while men slept, his enemy came and sowed

tares among the wheat, and went his way.

nes uenit inimicus eius et superseminauit zizania in medio tritici et abiit 13:26 cum autem creuisset herba et fructum 13:26 But when the blade was sprung up, and brought

forth fruit, then appeared the tares also.

fecisset tunc apparuerunt et zizania 13:27 accedentes autem 13:27 So the servants of the householder came and said

unto him, Sir, didst not thou sow good seed in thy field? from whence then hath it tares?

serui patris familias dixerunt ei domine nonne bonum semen seminasti in agro tuo unde ergo habet zizania 13:28 et ait il- 13:28 He said unto them, An enemy hath done this. The

servants said unto him, Wilt thou then that we go and gather them up?

lis inimicus homo hoc fecit serui autem dixerunt ei uis imus et colligimus ea 13:29 et ait non ne forte colligentes ziza- 13:29 But he said, Nay; lest while ye gather up the tares,

ye root up also the wheat with them.

nia eradicetis simul cum eis et triticum 13:30 sinite utraque 13:30 Let both grow together until the harvest: and in the

time of harvest I will say to the reapers, Gather ye together first the tares, and bind them in bundles to burn them: but

crescere usque ad messem et in tempore messis dicam mes- gather the wheat into my barn.

Biblia Sacra Vulgata

soribus colligite primum zizania et alligate ea fasciculos ad conburendum triticum autem congregate in horreum meum

13:31 Another parable put he forth unto them, saying, 13:31 aliam parabolam proposuit eis dicens simile est reg-

The kingdom of heaven is like to a grain of mustard seed, which a man took, and sowed in his field:

num caelorum grano sinapis quod accipiens homo seminauit

13:32 Which indeed is the least of all seeds: but when it is in agro suo 13:32 quod minimum quidem est omnibus segrown, it is the greatest among herbs, and becometh a tree, so that the birds of the air come and lodge in the branches

thereof. minibus cum autem creuerit maius est omnibus holeribus et

fit arbor ita ut uolucres caeli ueniant et habitent in ramis eius

13:33 Another parable spake he unto them; The kingdom 13:33 aliam parabolam locutus est eis simile est regnum cae-

of heaven is like unto leaven, which a woman took, and hid in three measures of meal, till the whole was leavened.

lorum fermento quod acceptum mulier abscondit in farinae

13:34 All these things spake Jesus unto the multitude in satis tribus donec fermentatum est totum 13:34 haec omparables; and without a parable spake he not unto them:

nia locutus est Iesus in parabolis ad turbas et sine parabolis

13:35 That it might be fulfilled which was spoken by the non loquebatur eis 13:35 ut impleretur quod dictum erat per

prophet, saying, I will open my mouth in parables; I will utter things which have been kept secret from the founda-

tion of the world. prophetam dicentem aperiam in parabolis os meum eructabo 13:36 Then Jesus sent the multitude away, and went into abscondita a constitutione mundi 13:36 tunc dimissis turbis

the house: and his disciples came unto him, saying, Declare unto us the parable of the tares of the field.

uenit in domum et accesserunt ad eum discipuli eius dicentes

13:37 He answered and said unto them, He that soweth the dissere nobis parabolam zizaniorum agri 13:37 qui respon-

good seed is the Son of man; 13:38 The field is the world; the good seed are the children dens ait qui seminat bonum semen est Filius hominis 13:38 of the kingdom; but the tares are the children of the wicked one;

ager autem est mundus bonum uero semen hii sunt filii regni

13:39 The enemy that sowed them is the devil; the harvest zizania autem filii sunt nequam 13:39 inimicus autem qui is the end of the world; and the reapers are the angels.

seminauit ea est diabolus messis uero consummatio saeculi

13:40 As therefore the tares are gathered and burned in the est messores autem angeli sunt 13:40 sicut ergo colliguntur

fire; so shall it be in the end of this world.

zizania et igni conburuntur sic erit in consummatione saeculi

13:41 The Son of man shall send forth his angels, and they 13:41 mittet Filius hominis angelos suos et colligent de regno

shall gather out of his kingdom all things that offend, and them which do iniquity;

eius omnia scandala et eos qui faciunt iniquitatem 13:42 et

13:42 And shall cast them into a furnace of fire: there shall be wailing and gnashing of teeth.

mittent eos in caminum ignis ibi erit fletus et stridor dentium

13:43 Then shall the righteous shine forth as the sun in the 13:43 tunc iusti fulgebunt sicut sol in regno Patris eorum qui

kingdom of their Father. Who hath ears to hear, let him hear.

habet aures audiat 13:44 simile est regnum caelorum the-

13:44 Again, the kingdom of heaven is like unto treasure hid in a field; the which when a man hath found, he hideth, and for joy thereof goeth and selleth all that he hath, and buyeth that field.

sauro abscondito in agro quem qui inuenit homo abscondit et prae gaudio illius uadit et uendit uniuersa quae habet et emit

13:45 Again, the kingdom of heaven is like unto a mer- agrum illum 13:45 iterum simile est regnum caelorum ho-

chant man, seeking goodly pearls:

13:46 Who, when he had found one pearl of great price, mini negotiatori quaerenti bonas margaritas 13:46 inuenta

went and sold all that he had, and bought it.

autem una pretiosa margarita abiit et uendidit omnia quae

13:47 Again, the kingdom of heaven is like unto a net, that habuit et emit eam 13:47 iterum simile est regnum caelowas cast into the sea, and gathered of every kind:

rum sagenae missae in mare et ex omni genere congreganti 13:48 quam cum impleta esset educentes et secus litus se- 13:48 Which, when it was full, they drew to shore, and sat

down, and gathered the good into vessels, but cast the bad away.

dentes elegerunt bonos in uasa malos autem foras miserunt 13:49 sic erit in consummatione saeculi exibunt angeli et se- 13:49 So shall it be at the end of the world: the angels shall

come forth, and sever the wicked from among the just,

parabunt malos de medio iustorum 13:50 et mittent eos in 13:50 And shall cast them into the furnace of fire: there

shall be wailing and gnashing of teeth.

caminum ignis ibi erit fletus et stridor dentium 13:51 intel- 13:51 Jesus saith unto them, Have ye understood all these

things? They say unto him, Yea, Lord.

lexistis haec omnia dicunt ei etiam 13:52 ait illis ideo omnis 13:52 Then said he unto them, Therefore every scribe

which is instructed unto the kingdom of heaven is like unto a man that is an householder, which bringeth forth out of

scriba doctus in regno caelorum similis est homini patri fa- his treasure things new and old. milias qui profert de thesauro suo noua et uetera 13:53 et 13:53 And it came to pass, that when Jesus had finished

these parables, he departed thence.

factum est cum consummasset Iesus parabolas istas transiit inde 13:54 et ueniens in patriam suam docebat eos in syna- 13:54 And when he was come into his own country, he

taught them in their synagogue, insomuch that they were astonished, and said, Whence hath this man this wisdom,

gogis eorum ita ut mirarentur et dicerent unde huic sapientia and these mighty works? haec et uirtutes 13:55 nonne hic est fabri filius nonne mater 13:55 Is not this the carpenter’s son? is not his mother

called Mary? and his brethren, James, and Joses, and Simon, and Judas?

eius dicitur Maria et fratres eius Iacobus et Ioseph et Simon et Iudas 13:56 et sorores eius nonne omnes apud nos sunt 13:56 And his sisters, are they not all with us? Whence

then hath this man all these things?

unde ergo huic omnia ista 13:57 et scandalizabantur in eo 13:57 And they were offended in him. But Jesus said unto

them, A prophet is not without honour, save in his own country, and in his own house.

Iesus autem dixit eis non est propheta sine honore nisi in patria sua et in domo sua 13:58 et non fecit ibi uirtutes multas 13:58 And he did not many mighty works there because

of their unbelief.

propter incredulitatem illorum

14:1 in illo tempore audiit Herodes tetrarcha famam Iesu 14:1 At that time Herod the tetrarch heard of the fame of

Jesus,

14:2 et ait pueris suis hic est Iohannes Baptista ipse surrexit 14:2 And said unto his servants, This is John the Baptist;

he is risen from the dead; and therefore mighty works do shew forth themselves in him.

a mortuis et ideo uirtutes inoperantur in eo 14:3 Herodes

14:3 For Herod had laid hold on John, and bound him, and put him in prison for Herodias’ sake, his brother Philip’s wife.

enim tenuit Iohannem et alligauit eum et posuit in carcere propter Herodiadem uxorem fratris sui 14:4 dicebat enim 14:4 For John said unto him, It is not lawful for thee to

have her.

illi Iohannes non licet tibi habere eam 14:5 et uolens illum 14:5 And when he would have put him to death, he feared

the multitude, because they counted him as a prophet.

occidere timuit populum quia sicut prophetam eum habebant 14:6 die autem natalis Herodis saltauit filia Herodiadis in me- 14:6 But when Herod’s birthday was kept, the daughter of

Herodias danced before them, and pleased Herod.

dio et placuit Herodi 14:7 unde cum iuramento pollicitus 14:7 Whereupon he promised with an oath to give her

whatsoever she would ask.

est ei dare quodcumque postulasset ab eo 14:8 at illa prae- 14:8 And she, being before instructed of her mother, said,

Give me here John Baptist’s head in a charger.

monita a matre sua da mihi inquit hic in disco caput Iohannis Baptistae 14:9 et contristatus est rex propter iuramentum 14:9 And the king was sorry: nevertheless for the oath’s

sake, and them which sat with him at meat, he commanded it to be given her.

autem et eos qui pariter recumbebant iussit dari 14:10 mi-

14:10 And he sent, and beheaded John in the prison.

sitque et decollauit Iohannem in carcere 14:11 et adlatum 14:11 And his head was brought in a charger, and given to

the damsel: and she brought it to her mother.

Biblia Sacra Vulgata

est caput eius in disco et datum est puellae et tulit matri suae

14:12 And his disciples came, and took up the body, and 14:12 et accedentes discipuli eius tulerunt corpus et sepelie-

buried it, and went and told Jesus.

14:13 When Jesus heard of it, he departed thence by ship runt illud et uenientes nuntiauerunt Iesu 14:13 quod cum

into a desert place apart: and when the people had heard thereof, they followed him on foot out of the cities.

audisset Iesus secessit inde in nauicula in locum desertum seorsum et cum audissent turbae secutae sunt eum pedestres

14:14 And Jesus went forth, and saw a great multitude, de ciuitatibus 14:14 et exiens uidit turbam multam et miser-

and was moved with compassion toward them, and he healed their sick.

tus est eius et curauit languidos eorum 14:15 uespere autem

14:15 And when it was evening, his disciples came to him, saying, This is a desert place, and the time is now past; send the multitude away, that they may go into the villages, and buy themselves victuals.

facto accesserunt ad eum discipuli eius dicentes desertus est locus et hora iam praeteriit dimitte turbas ut euntes in cas-

14:16 But Jesus said unto them, They need not depart; give tella emant sibi escas 14:16 Iesus autem dixit eis non habent

ye them to eat.

14:17 And they say unto him, We have here but five loaves, necesse ire date illis uos manducare 14:17 responderunt ei

and two fishes.

14:18 He said, Bring them hither to me. non habemus hic nisi quinque panes et duos pisces 14:18 14:19 And he commanded the multitude to sit down on qui ait eis adferte illos mihi huc 14:19 et cum iussisset tur-

the grass, and took the five loaves, and the two fishes, and looking up to heaven, he blessed, and brake, and gave the loaves to his disciples, and the disciples to the multitude. bam discumbere supra faenum acceptis quinque panibus et

duobus piscibus aspiciens in caelum benedixit et fregit et

14:20 And they did all eat, and were filled: and they took dedit discipulis panes discipuli autem turbis 14:20 et man-

up of the fragments that remained twelve baskets full.

ducauerunt omnes et saturati sunt et tulerunt reliquias duo-

14:21 And they that had eaten were about five thousand decim cofinos fragmentorum plenos 14:21 manducantium

men, beside women and children.

autem fuit numerus quinque milia uirorum exceptis mulie-

14:22 And straightway Jesus constrained his disciples to ribus et paruulis 14:22 et statim iussit discipulos ascendere

get into a ship, and to go before him unto the other side, while he sent the multitudes away.

in nauicula et praecedere eum trans fretum donec dimitte-

14:23 And when he had sent the multitudes away, he went ret turbas 14:23 et dimissa turba ascendit in montem solus

up into a mountain apart to pray: and when the evening was come, he was there alone.

orare uespere autem facto solus erat ibi 14:24 nauicula au-

14:24 But the ship was now in the midst of the sea, tossed with waves: for the wind was contrary.

tem in medio mari iactabatur fluctibus erat enim contrarius

14:25 And in the fourth watch of the night Jesus went unto uentus 14:25 quarta autem uigilia noctis uenit ad eos am-

them, walking on the sea. 14:26 And when the disciples saw him walking on the sea, bulans supra mare 14:26 et uidentes eum supra mare amthey were troubled, saying, It is a spirit; and they cried out for fear.

bulantem turbati sunt dicentes quia fantasma est et prae ti-

14:27 But straightway Jesus spake unto them, saying, Be more clamauerunt 14:27 statimque Iesus locutus est eis di-

of good cheer; it is I; be not afraid.

14:28 And Peter answered him and said, Lord, if it be cens habete fiduciam ego sum nolite timere 14:28 respon-

thou, bid me come unto thee on the water.

dens autem Petrus dixit Domine si tu es iube me uenire ad te

14:29 And he said, Come. And when Peter was come super aquas 14:29 at ipse ait ueni et descendens Petrus de

down out of the ship, he walked on the water, to go to Jesus.

nauicula ambulabat super aquam ut ueniret ad Iesum 14:30

14:30 But when he saw the wind boisterous, he was afraid; and beginning to sink, he cried, saying, Lord, save me.

uidens uero uentum ualidum timuit et cum coepisset mergi clamauit dicens Domine saluum me fac 14:31 et continuo 14:31 And immediately Jesus stretched forth his hand,

and caught him, and said unto him, O thou of little faith, wherefore didst thou doubt?

Iesus extendens manum adprehendit eum et ait illi modicae fidei quare dubitasti 14:32 et cum ascendissent in nauicu- 14:32 And when they were come into the ship, the wind

ceased.

lam cessauit uentus 14:33 qui autem in nauicula erant ue- 14:33 Then they that were in the ship came and wor-

shipped him, saying, Of a truth thou art the Son of God.

nerunt et adorauerunt eum dicentes uere Filius Dei es 14:34 14:34 And when they were gone over, they came into the

land of Gennesaret.

et cum transfretassent uenerunt in terram Gennesar 14:35 et 14:35 And when the men of that place had knowledge of

him, they sent out into all that country round about, and brought unto him all that were diseased;

cum cognouissent eum uiri loci illius miserunt in uniuersam regionem illam et obtulerunt ei omnes male habentes 14:36 14:36 And besought him that they might only touch the

hem of his garment: and as many as touched were made perfectly whole.

et rogabant eum ut uel fimbriam uestimenti eius tangerent et quicumque tetigerunt salui facti sunt

15:1 tunc accesserunt ad eum ab Hierosolymis scribae et 15:1 Then came to Jesus scribes and Pharisees, which

were of Jerusalem, saying,

Pharisaei dicentes 15:2 quare discipuli tui transgrediuntur 15:2 Why do thy disciples transgress the tradition of the

elders? for they wash not their hands when they eat bread.

traditionem seniorum non enim lauant manus suas cum panem manducant 15:3 ipse autem respondens ait illis quare 15:3 But he answered and said unto them, Why do ye also

transgress the commandment of God by your tradition?

et uos transgredimini mandatum Dei propter traditionem uestram 15:4 nam Deus dixit honora patrem et matrem et qui 15:4 For God commanded, saying, Honour thy father and

mother: and, He that curseth father or mother, let him die the death.

maledixerit patri uel matri morte moriatur 15:5 uos autem

15:5 But ye say, Whosoever shall say to his father or his mother, It is a gift, by whatsoever thou mightest be profited by me;

dicitis quicumque dixerit patri uel matri munus quodcumque est ex me tibi prodeeit 15:6 et non honorificabit pat- 15:6 And honour not his father or his mother, he shall be

free. Thus have ye made the commandment of God of none effect by your tradition.

rem suum aut matrem et irritum fecistis mandatum Dei propter traditionem uestram 15:7 hypocritae bene prophetauit 15:7 Ye hypocrites, well did Esaias prophesy of you, say-

ing,

de uobis Esaias dicens 15:8 populus hic labiis me hono- 15:8 This people draweth nigh unto me with their mouth,

and honoureth me with their lips; but their heart is far from me.

rat cor autem eorum longe est a me 15:9 sine causa autem

15:9 But in vain they do worship me, teaching for doctrines the commandments of men.

colunt me docentes doctrinas mandata hominum 15:10 et 15:10 And he called the multitude, and said unto them,

Hear, and understand:

conuocatis ad se turbis dixit eis audite et intellegite 15:11 15:11 Not that which goeth into the mouth defileth a man;

but that which cometh out of the mouth, this defileth a man.

non quod intrat in os coinquinat hominem sed quod procedit ex ore hoc coinquinat hominem 15:12 tunc accedentes 15:12 Then came his disciples, and said unto him, Know-

est thou that the Pharisees were offended, after they heard this saying?

discipuli eius dixerunt ei scis quia Pharisaei audito uerbo scandalizati sunt 15:13 at ille respondens ait omnis plan- 15:13 But he answered and said, Every plant, which my

heavenly Father hath not planted, shall be rooted up.

tatio quam non plantauit Pater meus caelestis eradicabitur 15:14 sinite illos caeci sunt duces caecorum caecus autem 15:14 Let them alone: they be blind leaders of the blind.

And if the blind lead the blind, both shall fall into the ditch.

si caeco ducatum praestet ambo in foueam cadunt 15:15 15:15 Then answered Peter and said unto him, Declare

unto us this parable.

respondens autem Petrus dixit ei edissere nobis parabolam

Biblia Sacra Vulgata

15:16 And Jesus said, Are ye also yet without understand- istam 15:16 at ille dixit adhuc et uos sine intellectu estis

ing?

15:17 Do not ye yet understand, that whatsoever entereth 15:17 non intellegitis quia omne quod in os intrat in uentrem

in at the mouth goeth into the belly, and is cast out into the draught?

uadit et in secessum emittitur 15:18 quae autem procedunt

15:18 But those things which proceed out of the mouth come forth from the heart; and they defile the man. 15:19 For out of the heart proceed evil thoughts, murders, de ore de corde exeunt et ea coinquinant hominem 15:19 adulteries, fornications, thefts, false witness, blasphemies:

de corde enim exeunt cogitationes malae homicidia adulte-

15:20 These are the things which defile a man: but to eat ria fornicationes furta falsa testimonia blasphemiae 15:20

with unwashen hands defileth not a man.

haec sunt quae coinquinant hominem non lotis autem mani-

15:21 Then Jesus went thence, and departed into the coasts bus manducare non coinquinat hominem 15:21 et egressus of Tyre and Sidon.

15:22 And, behold, a woman of Canaan came out of the inde Iesus secessit in partes Tyri et Sidonis 15:22 et ecce

same coasts, and cried unto him, saying, Have mercy on me, O Lord, thou son of David; my daughter is grievously

vexed with a devil. mulier chananea a finibus illis egressa clamauit dicens ei

miserere mei Domine Fili Dauid filia mea male a daemo-

15:23 But he answered her not a word. And his disciples nio uexatur 15:23 qui non respondit ei uerbum et acceden-

came and besought him, saying, Send her away; for she crieth after us.

tes discipuli eius rogabant eum dicentes dimitte eam quia

15:24 But he answered and said, I am not sent but unto the clamat post nos 15:24 ipse autem respondens ait non sum lost sheep of the house of Israel.

15:25 Then came she and worshipped him, saying, Lord, missus nisi ad oues quae perierunt domus Israhel 15:25 at

help me.

15:26 But he answered and said, It is not meet to take the illa uenit et adorauit eum dicens Domine adiuua me 15:26

children’s bread, and to cast it to dogs.

qui respondens ait non est bonum sumere panem filiorum et

15:27 And she said, Truth, Lord: yet the dogs eat of the mittere canibus 15:27 at illa dixit etiam Domine nam et ca-

crumbs which fall from their masters’ table.

telli edunt de micis quae cadunt de mensa dominorum suo-

15:28 Then Jesus answered and said unto her, O woman, rum 15:28 tunc respondens Iesus ait illi o mulier magna est

great is thy faith: be it unto thee even as thou wilt. And her daughter was made whole from that very hour.

fides tua fiat tibi sicut uis et sanata est filia illius ex illa hora

15:29 And Jesus departed from thence, and came nigh 15:29 et cum transisset inde Iesus uenit secus mare Galilaeae

unto the sea of Galilee; and went up into a mountain, and sat down there.

et ascendens in montem sedebat ibi 15:30 et accesserunt ad

15:30 And great multitudes came unto him, having with them those that were lame, blind, dumb, maimed, and many others, and cast them down at Jesus’ feet; and he healed them:

eum turbae multae habentes secum mutos clodos caecos debiles et alios multos et proiecerunt eos ad pedes eius et cu-

15:31 Insomuch that the multitude wondered, when they rauit eos 15:31 ita ut turbae mirarentur uidentes mutos lo-

saw the dumb to speak, the maimed to be whole, the lame to walk, and the blind to see: and they glorified the God of

Israel. quentes clodos ambulantes caecos uidentes et magnificabant 15:32 Then Jesus called his disciples unto him, and said, I Deum Israhel 15:32 Iesus autem conuocatis discipulis suis

have compassion on the multitude, because they continue with me now three days, and have nothing to eat: and I

will not send them away fasting, lest they faint in the way. dixit misereor turbae quia triduo iam perseuerant mecum et

non habent quod manducent et dimittere eos ieiunos nolo ne

15:33 And his disciples say unto him, Whence should we deficiant in uia 15:33 et dicunt ei discipuli unde ergo nobis

have so much bread in the wilderness, as to fill so great a multitude?

in deserto panes tantos ut saturemus turbam tantam 15:34

15:34 And Jesus saith unto them, How many loaves have ye? And they said, Seven, and a few little fishes.

et ait illis Iesus quot panes habetis at illi dixerunt septem et

Euangelium Secundum Matthaeum

paucos pisciculos 15:35 et praecepit turbae ut discumberet 15:35 And he commanded the multitude to sit down on the

ground.

super terram 15:36 et accipiens septem panes et pisces et 15:36 And he took the seven loaves and the fishes, and

gave thanks, and brake them, and gave to his disciples, and the disciples to the multitude.

gratias agens fregit et dedit discipulis suis et discipuli dederunt populo 15:37 et comederunt omnes et saturati sunt et 15:37 And they did all eat, and were filled: and they took

up of the broken meat that was left seven baskets full.

quod superfuit de fragmentis tulerunt septem sportas plenas 15:38 erant autem qui manducauerant quattuor milia homi- 15:38 And they that did eat were four thousand men, be-

side women and children.

num extra paruulos et mulieres 15:39 et dimissa turba as- 15:39 And he sent away the multitude, and took ship, and

came into the coasts of Magdala.

cendit in nauiculam et uenit in fines Magedan

16:1 et accesserunt ad eum Pharisaei et Sadducaei tem- 16:1 The Pharisees also with the Sadducees came, and

tempting desired him that he would shew them a sign from heaven.

ptantes et rogauerunt eum ut signum de caelo ostenderet eis 16:2 at ille respondens ait eis facto uespere dicitis sere- 16:2 He answered and said unto them, When it is evening,

ye say, It will be fair weather: for the sky is red.

num erit rubicundum est enim caelum 16:3 et mane hodie 16:3 And in the morning, It will be foul weather to day:

for the sky is red and lowering. O ye hypocrites, ye can discern the face of the sky; but can ye not discern the signs

tempestas rutilat enim triste caelum 16:4 faciem ergo caeli of the times?

16:4 A wicked and adulterous generation seeketh after a sign; and there shall no sign be given unto it, but the sign of the prophet Jonas. And he left them, and departed.

diiudicare nostis signa autem temporum non potestis generatio mala et adultera signum quaerit et signum non dabitur ei nisi signum Ionae et relictis illis abiit 16:5 et cum ue- 16:5 And when his disciples were come to the other side,

they had forgotten to take bread.

nissent discipuli eius trans fretum obliti sunt panes accipere 16:6 qui dixit illis intuemini et cauete a fermento Pharisaeo- 16:6 Then Jesus said unto them, Take heed and beware of

the leaven of the Pharisees and of the Sadducees.

rum et Sadducaeorum 16:7 at illi cogitabant inter se dicen- 16:7 And they reasoned among themselves, saying, It is

because we have taken no bread.

tes quia panes non accepimus 16:8 sciens autem Iesus dixit 16:8 Which when Jesus perceived, he said unto them, O ye

of little faith, why reason ye among yourselves, because ye have brought no bread?

quid cogitatis inter uos modicae fidei quia panes non habetis 16:9 nondum intellegitis neque recordamini quinque 16:9 Do ye not yet understand, neither remember the five

loaves of the five thousand, and how many baskets ye took up?

panum quinque milium hominum et quot cofinos sumpsis-

tis 16:10 neque septem panum quattuor milium hominum 16:10 Neither the seven loaves of the four thousand, and

how many baskets ye took up?

et quot sportas sumpsistis 16:11 quare non intellegitis quia 16:11 How is it that ye do not understand that I spake it

not to you concerning bread, that ye should beware of the leaven of the Pharisees and of the Sadducees?

non de pane dixi uobis cauete a fermento Pharisaeorum et Sadducaeorum 16:12 tunc intellexerunt quia non dixerit ca- 16:12 Then understood they how that he bade them not

beware of the leaven of bread, but of the doctrine of the Pharisees and of the Sadducees.

uendum a fermento panum sed a doctrina Pharisaeorum et Sadducaeorum 16:13 uenit autem Iesus in partes Caesareae 16:13 When Jesus came into the coasts of Caesarea

Philippi, he asked his disciples, saying, Whom do men say that I the Son of man am?

Philippi et interrogabat discipulos suos dicens quem dicunt homines esse Filium hominis 16:14 at illi dixerunt alii Io- 16:14 And they said, Some say that thou art John the

Baptist: some, Elias; and others, Jeremias, or one of the prophets.

hannem Baptistam alii autem Heliam alii uero Hieremiam aut unum ex prophetis 16:15 dicit illis uos autem quem me 16:15 He saith unto them, But whom say ye that I am?

Biblia Sacra Vulgata

16:16 And Simon Peter answered and said, Thou art the esse dicitis 16:16 respondens Simon Petrus dixit tu es Ch-

Christ, the Son of the living God.

16:17 And Jesus answered and said unto him, Blessed art ristus Filius Dei uiui 16:17 respondens autem Iesus dixit ei

thou, Simon Barjona: for flesh and blood hath not revealed it unto thee, but my Father which is in heaven.

beatus es Simon Bar Iona quia caro et sanguis non reuela-

16:18 And I say also unto thee, That thou art Peter, and uit tibi sed Pater meus qui in caelis est 16:18 et ego dico

upon this rock I will build my church; and the gates of hell shall not prevail against it.

tibi quia tu es Petrus et super hanc petram aedificabo ecclesiam meam et portae inferi non praeualebunt aduersum eam

16:19 And I will give unto thee the keys of the kingdom 16:19 et tibi dabo claues regni caelorum et quodcumque li-

of heaven: and whatsoever thou shalt bind on earth shall be bound in heaven: and whatsoever thou shalt loose on

earth shall be loosed in heaven. gaueris super terram erit ligatum in caelis et quodcumque 16:20 Then charged he his disciples that they should tell solueris super terram erit solutum in caelis 16:20 tunc prae-

no man that he was Jesus the Christ.

cepit discipulis suis ut nemini dicerent quia ipse esset Iesus

16:21 From that time forth began Jesus to shew unto his Christus 16:21 exinde coepit Iesus ostendere discipulis suis

disciples, how that he must go unto Jerusalem, and suffer many things of the elders and chief priests and scribes, and

be killed, and be raised again the third day. quia oporteret eum ire Hierosolymam et multa pati a senio-

ribus et scribis et principibus sacerdotum et occidi et tertia

16:22 Then Peter took him, and began to rebuke him, say- die resurgere 16:22 et adsumens eum Petrus coepit incre-

ing, Be it far from thee, Lord: this shall not be unto thee.

16:23 But he turned, and said unto Peter, Get thee behind pare illum dicens absit a te Domine non erit tibi hoc 16:23

me, Satan: thou art an offence unto me: for thou savourest not the things that be of God, but those that be of men.

qui conuersus dixit Petro uade post me Satana scandalum es mihi quia non sapis ea quae Dei sunt sed ea quae hominum

16:24 Then said Jesus unto his disciples, If any man will 16:24 tunc Iesus dixit discipulis suis si quis uult post me ue-

come after me, let him deny himself, and take up his cross, and follow me.

nire abneget semet ipsum et tollat crucem suam et sequatur

16:25 For whosoever will save his life shall lose it: and me 16:25 qui enim uoluerit animam suam saluam facere whosoever will lose his life for my sake shall find it.

perdet eam qui autem perdiderit animam suam propter me

16:26 For what is a man profited, if he shall gain the whole inueniet eam 16:26 quid enim prodeet homini si mundum world, and lose his own soul? or what shall a man give in exchange for his soul?

uniuersum lucretur animae uero suae detrimentum patiatur

16:27 For the Son of man shall come in the glory of his aut quam dabit homo commutationem pro anima sua 16:27

Father with his angels; and then he shall reward every man according to his works.

Filius enim hominis uenturus est in gloria Patris sui cum angelis suis et tunc reddet unicuique secundum opus eius

16:28 Verily I say unto you, There be some standing here, 16:28 amen dico uobis sunt quidam de hic stantibus qui non

which shall not taste of death, till they see the Son of man coming in his kingdom.

gustabunt mortem donec uideant Filium hominis uenientem in regno suo

17:1 And after six days Jesus taketh Peter, James, and John 17:1 et post dies sex adsumpsit Iesus Petrum et Iacobum his brother, and bringeth them up into an high mountain apart,

et Iohannem fratrem eius et ducit illos in montem excelsum

17:2 And was transfigured before them: and his face did seorsum 17:2 et transfiguratus est ante eos et resplenduit

shine as the sun, and his raiment was white as the light.

facies eius sicut sol uestimenta autem eius facta sunt alba

Euangelium Secundum Matthaeum

sicut nix 17:3 et ecce apparuit illis Moses et Helias cum 17:3 And, behold, there appeared unto them Moses and

Elias talking with him.

eo loquentes 17:4 respondens autem Petrus dixit ad Iesum 17:4 Then answered Peter, and said unto Jesus, Lord, it is

good for us to be here: if thou wilt, let us make here three tabernacles; one for thee, and one for Moses, and one for

Domine bonum est nos hic esse si uis faciamus hic tria ta- Elias.

bernacula tibi unum et Mosi unum et Heliae unum 17:5 ad- 17:5 While he yet spake, behold, a bright cloud overshadowed them: and behold a voice out of the cloud, which said, This is my beloved Son, in whom I am well pleased;

huc eo loquente ecce nubes lucida obumbrauit eos et ecce hear ye him. uox de nube dicens hic est Filius meus dilectus in quo mihi bene conplacuit ipsum audite 17:6 et audientes discipuli ce- 17:6 And when the disciples heard it, they fell on their

face, and were sore afraid.

ciderunt in faciem suam et timuerunt ualde 17:7 et acces- 17:7 And Jesus came and touched them, and said, Arise,

and be not afraid.

sit Iesus et tetigit eos dixitque eis surgite et nolite timere 17:8 leuantes autem oculos suos neminem uiderunt nisi so- 17:8 And when they had lifted up their eyes, they saw no

man, save Jesus only.

lum Iesum 17:9 et descendentibus illis de monte praecepit 17:9 And as they came down from the mountain, Jesus

charged them, saying, Tell the vision to no man, until the Son of man be risen again from the dead.

Iesus dicens nemini dixeritis uisionem donec Filius hominis a mortuis resurgat 17:10 et interrogauerunt eum discipuli 17:10 And his disciples asked him, saying, Why then say

the scribes that Elias must first come?

dicentes quid ergo scribae dicunt quod Heliam oporteat primum uenire 17:11 at ille respondens ait eis Helias quidem 17:11 And Jesus answered and said unto them, Elias truly

shall first come, and restore all things.

uenturus est et restituet omnia 17:12 dico autem uobis quia 17:12 But I say unto you, That Elias is come already, and

they knew him not, but have done unto him whatsoever they listed. Likewise shall also the Son of man suffer of

Helias iam uenit et non cognouerunt eum sed fecerunt in eo them. quaecumque uoluerunt sic et Filius hominis passurus est ab eis 17:13 tunc intellexerunt discipuli quia de Iohanne Bap- 17:13 Then the disciples understood that he spake unto

them of John the Baptist.

tista dixisset eis 17:14 et cum uenisset ad turbam accessit 17:14 And when they were come to the multitude, there

came to him a certain man, kneeling down to him, and saying, Lord, have mercy on my son: for he is lunatick, and sore vexed: for ofttimes he falleth into the fire, and oft into the water.

ad eum homo genibus prouolutus ante eum dicens Domine miserere filii mei quia lunaticus est et male patitur nam saepe cadit in ignem et crebro in aquam 17:15 et obtuli eum dis- 17:15 And I brought him to thy disciples, and they could

not cure him.

cipulis tuis et non potuerunt curare eum 17:16 respondens 17:16 Then Jesus answered and said, O faithless and per-

verse generation, how long shall I be with you? how long shall I suffer you? bring him hither to me.

Iesus ait o generatio incredula et peruersa quousque ero uobiscum usquequo patiar uos adferte huc illum ad me 17:17 17:17 And Jesus rebuked the devil; and he departed out of

him: and the child was cured from that very hour.

et increpauit ei Iesus et exiit ab eo daemonium et curatus est puer ex illa hora 17:18 tunc accesserunt discipuli ad Ie- 17:18 Then came the disciples to Jesus apart, and said,

Why could not we cast him out?

sum secreto et dixerunt quare nos non potuimus eicere illum 17:19 dicit illis propter incredulitatem uestram amen quippe 17:19 And Jesus said unto them, Because of your unbelief:

for verily I say unto you, If ye have faith as a grain of mustard seed, ye shall say unto this mountain, Remove hence to yonder place; and it shall remove; and nothing shall be impossible unto you.

dico uobis si habueritis fidem sicut granum sinapis dicetis monti huic transi hinc et transibit et nihil inpossibile erit uobis 17:20 hoc autem genus non eicitur nisi per orationem et 17:20 Howbeit this kind goeth not out but by prayer and

fasting.

Biblia Sacra Vulgata

17:21 And while they abode in Galilee, Jesus said unto ieiunium 17:21 conuersantibus autem eis in Galilaea dixit

them, The Son of man shall be betrayed into the hands of men:

illis Iesus Filius hominis tradendus est in manus hominum

17:22 And they shall kill him, and the third day he shall 17:22 et occident eum et tertio die resurget et contristati sunt

be raised again. And they were exceeding sorry.

17:23 And when they were come to Capernaum, they that uehementer 17:23 et cum uenissent Capharnaum accesse-

received tribute money came to Peter, and said, Doth not your master pay tribute?

runt qui didragma accipiebant ad Petrum et dixerunt magi-

17:24 He saith, Yes. And when he was come into the ster uester non soluit didragma 17:24 ait etiam et cum in-

house, Jesus prevented him, saying, What thinkest thou, Simon? of whom do the kings of the earth take custom or

tribute? of their own children, or of strangers? trasset domum praeuenit eum Iesus dicens quid tibi uidetur

Simon reges terrae a quibus accipiunt tributum uel censum

17:25 Peter saith unto him, Of strangers. Jesus saith unto a filiis suis an ab alienis 17:25 et ille dixit ab alienis dixit

him, Then are the children free.

17:26 Notwithstanding, lest we should offend them, go illi Iesus ergo liberi sunt filii 17:26 ut autem non scandali-

thou to the sea, and cast an hook, and take up the fish that first cometh up; and when thou hast opened his mouth, thou shalt find a piece of money: that take, and give unto them for me and thee.

zemus eos uade ad mare et mitte hamum et eum piscem qui primus ascenderit tolle et aperto ore eius inuenies staterem illum sumens da eis pro me et te

18:1 At the same time came the disciples unto Jesus, say- 18:1 in illa hora accesserunt discipuli ad Iesum dicentes

ing, Who is the greatest in the kingdom of heaven? 18:2 And Jesus called a little child unto him, and set him quis putas maior est in regno caelorum 18:2 et aduocans in the midst of them,

18:3 And said, Verily I say unto you, Except ye be con- Iesus paruulum statuit eum in medio eorum 18:3 et dixit

verted, and become as little children, ye shall not enter into the kingdom of heaven.

amen dico uobis nisi conuersi fueritis et efficiamini sicut

18:4 Whosoever therefore shall humble himself as this lit- paruuli non intrabitis in regnum caelorum 18:4 quicumque tle child, the same is greatest in the kingdom of heaven.

ergo humiliauerit se sicut paruulus iste hic est maior in re-

18:5 And whoso shall receive one such little child in my gno caelorum 18:5 et qui susceperit unum paruulum talem

name receiveth me.

18:6 But whoso shall offend one of these little ones which in nomine meo me suscipit 18:6 qui autem scandalizauerit

believe in me, it were better for him that a millstone were hanged about his neck, and that he were drowned in the

depth of the sea. unum de pusillis istis qui in me credunt expedit ei ut suspen-

datur mola asinaria in collo eius et demergatur in profundum

18:7 Woe unto the world because of offences! for it must maris 18:7 uae mundo ab scandalis necesse est enim ut ue-

needs be that offences come; but woe to that man by whom the offence cometh!

niant scandala uerumtamen uae homini per quem scanda-

18:8 Wherefore if thy hand or thy foot offend thee, cut lum uenit 18:8 si autem manus tua uel pes tuus scandalizat

them off, and cast them from thee: it is better for thee to enter into life halt or maimed, rather than having two

hands or two feet to be cast into everlasting fire. te abscide eum et proice abs te bonum tibi est ad uitam in-

gredi debilem uel clodum quam duas manus uel duos pedes

18:9 And if thine eye offend thee, pluck it out, and cast it habentem mitti in ignem aeternum 18:9 et si oculus tuus from thee: it is better for thee to enter into life with one eye, rather than having two eyes to be cast into hell fire.

scandalizat te erue eum et proice abs te bonum tibi est unoculum in uitam intrare quam duos oculos habentem mitti in

18:10 Take heed that ye despise not one of these little ones; gehennam ignis 18:10 uidete ne contemnatis unum ex his for I say unto you, That in heaven their angels do always behold the face of my Father which is in heaven.

pusillis dico enim uobis quia angeli eorum in caelis sem-

Euangelium Secundum Matthaeum

per uident faciem Patris mei qui in caelis est 18:11 uenit 18:11 For the Son of man is come to save that which was

lost.

enim Filius hominis saluare quod perierat 18:12 quid uo- 18:12 How think ye? if a man have an hundred sheep, and

one of them be gone astray, doth he not leave the ninety and nine, and goeth into the mountains, and seeketh that

bis uidetur si fuerint alicui centum oues et errauerit una ex which is gone astray? eis nonne relinquet nonaginta nouem in montibus et uadit quaerere eam quae errauit 18:13 et si contigerit ut inueniat 18:13 And if so be that he find it, verily I say unto you, he

rejoiceth more of that sheep, than of the ninety and nine which went not astray.

eam amen dico uobis quia gaudebit super eam magis quam super nonaginta nouem quae non errauerunt 18:14 sic non 18:14 Even so it is not the will of your Father which is in

heaven, that one of these little ones should perish.

est uoluntas ante Patrem uestrum qui in caelis est ut pereat unus de pusillis istis 18:15 si autem peccauerit in te frater 18:15 Moreover if thy brother shall trespass against thee,

go and tell him his fault between thee and him alone: if he shall hear thee, thou hast gained thy brother.

tuus uade et corripe eum inter te et ipsum solum si te audierit lucratus es fratrem tuum 18:16 si autem non te audierit 18:16 But if he will not hear thee, then take with thee one

or two more, that in the mouth of two or three witnesses every word may be established.

adhibe tecum adhuc unum uel duos ut in ore duorum testium uel trium stet omne uerbum 18:17 quod si non audierit 18:17 And if he shall neglect to hear them, tell it unto the

church: but if he neglect to hear the church, let him be unto thee as an heathen man and a publican.

eos dic ecclesiae si autem et ecclesiam non audierit sit tibi sicut ethnicus et publicanus 18:18 amen dico uobis quae- 18:18 Verily I say unto you, Whatsoever ye shall bind on

earth shall be bound in heaven: and whatsoever ye shall loose on earth shall be loosed in heaven.

cumque alligaueritis super terram erunt ligata et in caelo et quaecumque solueritis super terram erunt soluta et in caelo 18:19 iterum dico uobis quia si duo ex uobis consenserint su- 18:19 Again I say unto you, That if two of you shall agree

on earth as touching any thing that they shall ask, it shall be done for them of my Father which is in heaven.

per terram de omni re quacumque petierint fiet illis a Patre meo qui in caelis est 18:20 ubi enim sunt duo uel tres con- 18:20 For where two or three are gathered together in my

name, there am I in the midst of them.

gregati in nomine meo ibi sum in medio eorum 18:21 tunc 18:21 Then came Peter to him, and said, Lord, how oft

shall my brother sin against me, and I forgive him? till seven times?

accedens Petrus ad eum dixit Domine quotiens peccabit in me frater meus et dimittam ei usque septies 18:22 dicit illi 18:22 Jesus saith unto him, I say not unto thee, Until seven

times: but, Until seventy times seven.

Iesus non dico tibi usque septies sed usque septuagies septies 18:23 ideo adsimilatum est regnum caelorum homini regi qui 18:23 Therefore is the kingdom of heaven likened unto a

certain king, which would take account of his servants.

uoluit rationem ponere cum seruis suis 18:24 et cum coe- 18:24 And when he had begun to reckon, one was brought

unto him, which owed him ten thousand talents.

pisset rationem ponere oblatus est ei unus qui debebat decem milia talenta 18:25 cum autem non haberet unde red- 18:25 But forasmuch as he had not to pay, his lord com-

manded him to be sold, and his wife, and children, and all that he had, and payment to be made.

deret iussit eum dominus uenundari et uxorem eius et filios et omnia quae habebat et reddi 18:26 procidens autem ser- 18:26 The servant therefore fell down, and worshipped

him, saying, Lord, have patience with me, and I will pay thee all.

uus ille orabat eum dicens patientiam habe in me et omnia reddam tibi 18:27 misertus autem dominus serui illius dimi- 18:27 Then the lord of that servant was moved with com-

passion, and loosed him, and forgave him the debt.

sit eum et debitum dimisit ei 18:28 egressus autem seruus 18:28 But the same servant went out, and found one of his

fellowservants, which owed him an hundred pence: and he laid hands on him, and took him by the throat, saying, Pay me that thou owest.

Biblia Sacra Vulgata

ille inuenit unum de conseruis suis qui debebat ei centum denarios et tenens suffocabat eum dicens redde quod debes

18:29 And his fellowservant fell down at his feet, and be- 18:29 et procidens conseruus eius rogabat eum dicens patien-

sought him, saying, Have patience with me, and I will pay thee all.

tiam habe in me et omnia reddam tibi 18:30 ille autem no-

18:30 And he would not: but went and cast him into prison, till he should pay the debt.

luit sed abiit et misit eum in carcerem donec redderet debi-

18:31 So when his fellowservants saw what was done, they tum 18:31 uidentes autem conserui eius quae fiebant contri-

were very sorry, and came and told unto their lord all that was done.

stati sunt ualde et uenerunt et narrauerunt domino suo omnia

18:32 Then his lord, after that he had called him, said unto quae facta erant 18:32 tunc uocauit illum dominus suus et him, O thou wicked servant, I forgave thee all that debt, because thou desiredst me:

ait illi serue nequam omne debitum dimisi tibi quoniam ro-

18:33 Shouldest not thou also have had compassion on thy gasti me 18:33 non ergo oportuit et te misereri conserui tui

fellowservant, even as I had pity on thee?

18:34 And his lord was wroth, and delivered him to the sicut et ego tui misertus sum 18:34 et iratus dominus eius

tormentors, till he should pay all that was due unto him.

tradidit eum tortoribus quoadusque redderet uniuersum de-

18:35 So likewise shall my heavenly Father do also unto bitum 18:35 sic et Pater meus caelestis faciet uobis si non

you, if ye from your hearts forgive not every one his brother their trespasses.

remiseritis unusquisque fratri suo de cordibus uestris

19:1 And it came to pass, that when Jesus had finished 19:1 et factum est cum consummasset Iesus sermones is-

these sayings, he departed from Galilee, and came into the coasts of Judaea beyond Jordan;

tos migrauit a Galilaea et uenit in fines Iudaeae trans Iorda-

19:2 And great multitudes followed him; and he healed nen 19:2 et secutae sunt eum turbae multae et curauit eos

them there. 19:3 The Pharisees also came unto him, tempting him, and ibi 19:3 et accesserunt ad eum Pharisaei temptantes eum saying unto him, Is it lawful for a man to put away his wife for every cause?

et dicentes si licet homini dimittere uxorem suam quacum-

19:4 And he answered and said unto them, Have ye not que ex causa 19:4 qui respondens ait eis non legistis quia

read, that he which made them at the beginning made them male and female,

qui fecit ab initio masculum et feminam fecit eos 19:5 et

19:5 And said, For this cause shall a man leave father and mother, and shall cleave to his wife: and they twain shall be one flesh?

dixit propter hoc dimittet homo patrem et matrem et adhe-

19:6 Wherefore they are no more twain, but one flesh. rebit uxori suae et erunt duo in carne una 19:6 itaque iam

What therefore God hath joined together, let not man put asunder.

non sunt duo sed una caro quod ergo Deus coniunxit homo

19:7 They say unto him, Why did Moses then command non separet 19:7 dicunt illi quid ergo Moses mandauit dari

to give a writing of divorcement, and to put her away?

19:8 He saith unto them, Moses because of the hardness libellum repudii et dimittere 19:8 ait illis quoniam Moses

of your hearts suffered you to put away your wives: but from the beginning it was not so.

ad duritiam cordis uestri permisit uobis dimittere uxores ue-

19:9 And I say unto you, Whosoever shall put away his stras ab initio autem non sic fuit 19:9 dico autem uobis quia

wife, except it be for fornication, and shall marry another, committeth adultery: and whoso marrieth her which is put

away doth commit adultery. quicumque dimiserit uxorem suam nisi ob fornicationem et

aliam duxerit moechatur et qui dimissam duxerit moechatur

19:10 His disciples say unto him, If the case of the man be 19:10 dicunt ei discipuli eius si ita est causa homini cum ux-

so with his wife, it is not good to marry.

19:11 But he said unto them, All men cannot receive this ore non expedit nubere 19:11 qui dixit non omnes capiunt

saying, save they to whom it is given.

19:12 For there are some eunuchs, which were so born uerbum istud sed quibus datum est 19:12 sunt enim eunuchi

from their mother’s womb: and there are some eunuchs, which were made eunuchs of men: and there be eunuchs, which have made themselves eunuchs for the kingdom of heaven’s sake. He that is able to receive it, let him receive it.

MCCCLXXVI

Euangelium Secundum Matthaeum

qui de matris utero sic nati sunt et sunt eunuchi qui facti sunt ab hominibus et sunt eunuchi qui se ipsos castrauerunt propter regnum caelorum qui potest capere capiat 19:13 tunc 19:13 Then were there brought unto him little children,

that he should put his hands on them, and pray: and the disciples rebuked them.

oblati sunt ei paruuli ut manus eis inponeret et oraret discipuli autem increpabant eis 19:14 Iesus uero ait eis si- 19:14 But Jesus said, Suffer little children, and forbid

them not, to come unto me: for of such is the kingdom of heaven.

nite paruulos et nolite eos prohibere ad me uenire talium est

enim regnum caelorum 19:15 et cum inposuisset eis ma- 19:15 And he laid his hands on them, and departed thence.

nus abiit inde 19:16 et ecce unus accedens ait illi magister 19:16 And, behold, one came and said unto him, Good

Master, what good thing shall I do, that I may have eternal life?

bone quid boni faciam ut habeam uitam aeternam 19:17 qui

19:17 And he said unto him, Why callest thou me good? there is none good but one, that is, God: but if thou wilt enter into life, keep the commandments.

dixit ei quid me interrogas de bono unus est bonus Deus si autem uis ad uitam ingredi serua mandata 19:18 dicit illi 19:18 He saith unto him, Which? Jesus said, Thou shalt

do no murder, Thou shalt not commit adultery, Thou shalt not steal, Thou shalt not bear false witness,

quae Iesus autem dixit non homicidium facies non adulterabis non facies furtum non falsum testimonium dices 19:19 19:19 Honour thy father and thy mother: and, Thou shalt

love thy neighbour as thyself.

honora patrem et matrem et diliges proximum tuum sicut te ipsum 19:20 dicit illi adulescens omnia haec custodiui quid 19:20 The young man saith unto him, All these things have

I kept from my youth up: what lack I yet?

adhuc mihi deest 19:21 ait illi Iesus si uis perfectus esse 19:21 Jesus said unto him, If thou wilt be perfect, go and

sell that thou hast, and give to the poor, and thou shalt have treasure in heaven: and come and follow me.

uade uende quae habes et da pauperibus et habebis thesaurum in caelo et ueni sequere me 19:22 cum audisset autem 19:22 But when the young man heard that saying, he went

away sorrowful: for he had great possessions.

adulescens uerbum abiit tristis erat enim habens multas possessiones 19:23 Iesus autem dixit discipulis suis amen dico 19:23 Then said Jesus unto his disciples, Verily I say unto

you, That a rich man shall hardly enter into the kingdom of heaven.

uobis quia diues difficile intrabit in regnum caelorum 19:24

19:24 And again I say unto you, It is easier for a camel to go through the eye of a needle, than for a rich man to enter into the kingdom of God.

et iterum dico uobis facilius est camelum per foramen acus transire quam diuitem intrare in regnum caelorum 19:25 au- 19:25 When his disciples heard it, they were exceedingly

amazed, saying, Who then can be saved?

ditis autem his discipuli mirabantur ualde dicentes quis ergo poterit saluus esse 19:26 aspiciens autem Iesus dixit illis 19:26 But Jesus beheld them, and said unto them, With

men this is impossible; but with God all things are possible.

apud homines hoc inpossibile est apud Deum autem omnia possibilia sunt 19:27 tunc respondens Petrus dixit ei ecce 19:27 Then answered Peter and said unto him, Behold, we

have forsaken all, and followed thee; what shall we have therefore?

nos reliquimus omnia et secuti sumus te quid ergo erit nobis 19:28 Iesus autem dixit illis amen dico uobis quod uos qui se- 19:28 And Jesus said unto them, Verily I say unto you,

That ye which have followed me, in the regeneration when the Son of man shall sit in the throne of his glory, ye also shall sit upon twelve thrones, judging the twelve tribes of Israel.

cuti estis me in regeneratione cum sederit Filius hominis in sede maiestatis suae sedebitis et uos super sedes duodecim iudicantes duodecim tribus Israhel 19:29 et omnis qui reli- 19:29 And every one that hath forsaken houses, or

brethren, or sisters, or father, or mother, or wife, or children, or lands, for my name’s sake, shall receive an hun-

quit domum uel fratres aut sorores aut patrem aut matrem dredfold, and shall inherit everlasting life.

Biblia Sacra Vulgata

aut uxorem aut filios aut agros propter nomen meum cen-

19:30 But many that are first shall be last; and the last shall tuplum accipiet et uitam aeternam possidebit 19:30 multi be first.

autem erunt primi nouissimi et nouissimi primi

20:1 For the kingdom of heaven is like unto a man that is 20:1 simile est enim regnum caelorum homini patri faan householder, which went out early in the morning to hire labourers into his vineyard.

milias qui exiit primo mane conducere operarios in uineam

20:2 And when he had agreed with the labourers for a suam 20:2 conuentione autem facta cum operariis ex dena-

penny a day, he sent them into his vineyard.

20:3 And he went out about the third hour, and saw others rio diurno misit eos in uineam suam 20:3 et egressus circa

standing idle in the marketplace,

20:4 And said unto them; Go ye also into the vineyard, horam tertiam uidit alios stantes in foro otiosos 20:4 et illis

and whatsoever is right I will give you. And they went their way.

dixit ite et uos in uineam et quod iustum fuerit dabo uobis

20:5 Again he went out about the sixth and ninth hour, and 20:5 illi autem abierunt iterum autem exiit circa sextam et

did likewise.

20:6 And about the eleventh hour he went out, and found nonam horam et fecit similiter 20:6 circa undecimam uero

others standing idle, and saith unto them, Why stand ye here all the day idle?

exiit et inuenit alios stantes et dicit illis quid hic statis tota

20:7 They say unto him, Because no man hath hired us. He die otiosi 20:7 dicunt ei quia nemo nos conduxit dicit illis

saith unto them, Go ye also into the vineyard; and whatsoever is right, that shall ye receive.

ite et uos in uineam 20:8 cum sero autem factum esset dicit

20:8 So when even was come, the lord of the vineyard saith unto his steward, Call the labourers, and give them their hire, beginning from the last unto the first.

dominus uineae procuratori suo uoca operarios et redde il-

20:9 And when they came that were hired about the lis mercedem incipiens a nouissimis usque ad primos 20:9

eleventh hour, they received every man a penny.

cum uenissent ergo qui circa undecimam horam uenerant ac-

20:10 But when the first came, they supposed that they ceperunt singulos denarios 20:10 uenientes autem et primi

should have received more; and they likewise received every man a penny.

arbitrati sunt quod plus essent accepturi acceperunt autem

20:11 And when they had received it, they murmured et ipsi singulos denarios 20:11 et accipientes murmurabant

against the goodman of the house,

20:12 Saying, These last have wrought but one hour, and aduersus patrem familias 20:12 dicentes hii nouissimi una

thou hast made them equal unto us, which have borne the burden and heat of the day.

hora fecerunt et pares illos nobis fecisti qui portauimus pon-

20:13 But he answered one of them, and said, Friend, I do dus diei et aestus 20:13 at ille respondens uni eorum dixit

thee no wrong: didst not thou agree with me for a penny?

amice non facio tibi iniuriam nonne ex denario conuenisti

20:14 Take that thine is, and go thy way: I will give unto mecum 20:14 tolle quod tuum est et uade uolo autem et this last, even as unto thee.

20:15 Is it not lawful for me to do what I will with mine huic nouissimo dare sicut et tibi 20:15 aut non licet mihi

own? Is thine eye evil, because I am good?

quod uolo facere an oculus tuus nequam est quia ego bo-

20:16 So the last shall be first, and the first last: for many nus sum 20:16 sic erunt nouissimi primi et primi nouissimi

be called, but few chosen.

20:17 And Jesus going up to Jerusalem took the twelve multi sunt enim uocati pauci autem electi 20:17 et ascen-

disciples apart in the way, and said unto them,

dens Iesus Hierosolymam adsumpsit duodecim discipulos

20:18 Behold, we go up to Jerusalem; and the Son of secreto et ait illis 20:18 ecce ascendimus Hierosolymam et

man shall be betrayed unto the chief priests and unto the scribes, and they shall condemn him to death,

Filius hominis tradetur principibus sacerdotum et scribis et

20:19 And shall deliver him to the Gentiles to mock, and condemnabunt eum morte 20:19 et tradent eum gentibus ad

to scourge, and to crucify him: and the third day he shall rise again.

Euangelium Secundum Matthaeum

deludendum et flagellandum et crucifigendum et tertia die resurget 20:20 tunc accessit ad eum mater filiorum Zebe- 20:20 Then came to him the mother of Zebedees children

with her sons, worshipping him, and desiring a certain thing of him.

daei cum filiis suis adorans et petens aliquid ab eo 20:21 qui

20:21 And he said unto her, What wilt thou? She saith unto him, Grant that these my two sons may sit, the one on thy right hand, and the other on the left, in thy kingdom.

dixit ei quid uis ait illi dic ut sedeant hii duo filii mei unus ad dexteram tuam et unus ad sinistram in regno tuo 20:22 20:22 But Jesus answered and said, Ye know not what ye

ask. Are ye able to drink of the cup that I shall drink of, and to be baptized with the baptism that I am baptized

respondens autem Iesus dixit nescitis quid petatis potestis with? They say unto him, We are able. bibere calicem quem ego bibiturus sum dicunt ei possumus 20:23 ait illis calicem quidem meum bibetis sedere autem ad 20:23 And he saith unto them, Ye shall drink indeed of my

cup, and be baptized with the baptism that I am baptized with: but to sit on my right hand, and on my left, is not mine to give, but it shall be given to them for whom it is prepared of my Father.

dexteram meam et sinistram non est meum dare uobis sed quibus paratum est a Patre meo 20:24 et audientes decem 20:24 And when the ten heard it, they were moved with

indignation against the two brethren.

indignati sunt de duobus fratribus 20:25 Iesus autem uoca- 20:25 But Jesus called them unto him, and said, Ye know

that the princes of the Gentiles exercise dominion over them, and they that are great exercise authority upon them.

uit eos ad se et ait scitis quia principes gentium dominantur eorum et qui maiores sunt potestatem exercent in eos 20:26 20:26 But it shall not be so among you: but whosoever

will be great among you, let him be your minister;

non ita erit inter uos sed quicumque uoluerit inter uos maior fieri sit uester minister 20:27 et qui uoluerit inter uos pri- 20:27 And whosoever will be chief among you, let him be

your servant:

mus esse erit uester seruus 20:28 sicut Filius hominis non 20:28 Even as the Son of man came not to be ministered

unto, but to minister, and to give his life a ransom for many.

uenit ministrari sed ministrare et dare animam suam redeeptionem pro multis 20:29 et egredientibus eis ab Hiericho 20:29 And as they departed from Jericho, a great multitude

followed him.

secuta est eum turba multa 20:30 et ecce duo caeci seden- 20:30 And, behold, two blind men sitting by the way side,

when they heard that Jesus passed by, cried out, saying, Have mercy on us, O Lord, thou son of David.

tes secus uiam audierunt quia Iesus transiret et clamauerunt dicentes Domine miserere nostri Fili Dauid 20:31 turba au- 20:31 And the multitude rebuked them, because they

should hold their peace: but they cried the more, saying, Have mercy on us, O Lord, thou son of David.

tem increpabat eos ut tacerent at illi magis clamabant dicentes Domine miserere nostri Fili Dauid 20:32 et stetit Iesus 20:32 And Jesus stood still, and called them, and said,

What will ye that I shall do unto you?

et uocauit eos et ait quid uultis ut faciam uobis 20:33 dicunt 20:33 They say unto him, Lord, that our eyes may be

opened.

illi Domine ut aperiantur oculi nostri 20:34 misertus autem 20:34 So Jesus had compassion on them, and touched their

eyes: and immediately their eyes received sight, and they followed him.

eorum Iesus tetigit oculos eorum et confestim uiderunt et secuti sunt eum

21:1 et cum adpropinquassent Hierosolymis et uenissent 21:1 And when they drew nigh unto Jerusalem, and were

come to Bethphage, unto the mount of Olives, then sent Jesus two disciples,

Bethfage ad montem Oliueti tunc Iesus misit duos discipulos 21:2 dicens eis ite in castellum quod contra uos est et 21:2 Saying unto them, Go into the village over against

you, and straightway ye shall find an ass tied, and a colt with her: loose them, and bring them unto me.

statim inuenietis asinam alligatam et pullum cum ea soluite et adducite mihi 21:3 et si quis uobis aliquid dixerit dicite 21:3 And if any man say ought unto you, ye shall say,

The Lord hath need of them; and straightway he will send them.

quia Dominus his opus habet et confestim dimittet eos 21:4

21:4 All this was done, that it might be fulfilled which was spoken by the prophet, saying,

Biblia Sacra Vulgata

hoc autem factum est ut impleretur quod dictum est per pro-

21:5 Tell ye the daughter of Sion, Behold, thy King phetam dicentem 21:5 dicite filiae Sion ecce rex tuus uenit

cometh unto thee, meek, and sitting upon an ass, and a colt the foal of an ass.

tibi mansuetus et sedens super asinam et pullum filium su-

21:6 And the disciples went, and did as Jesus commanded biugalis 21:6 euntes autem discipuli fecerunt sicut praece-

them,

21:7 And brought the ass, and the colt, and put on them pit illis Iesus 21:7 et adduxerunt asinam et pullum et inpo-

their clothes, and they set him thereon.

suerunt super eis uestimenta sua et eum desuper sedere fe-

21:8 And a very great multitude spread their garments in cerunt 21:8 plurima autem turba strauerunt uestimenta sua

the way; others cut down branches from the trees, and strawed them in the way.

in uia alii autem caedebant ramos de arboribus et sternebant

21:9 And the multitudes that went before, and that fol- in uia 21:9 turbae autem quae praecedebant et quae seque-

lowed, cried, saying, Hosanna to the son of David: Blessed is he that cometh in the name of the Lord;

Hosanna in the highest. bantur clamabant dicentes osanna Filio Dauid benedictus qui 21:10 And when he was come into Jerusalem, all the city uenturus est in nomine Domini osanna in altissimis 21:10

was moved, saying, Who is this?

et cum intrasset Hierosolymam commota est uniuersa ciui-

21:11 And the multitude said, This is Jesus the prophet of tas dicens quis est hic 21:11 populi autem dicebant hic est

Nazareth of Galilee.

21:12 And Jesus went into the temple of God, and cast out Iesus propheta a Nazareth Galilaeae 21:12 et intrauit Iesus

all them that sold and bought in the temple, and overthrew the tables of the moneychangers, and the seats of them that

sold doves, in templum Dei et eiciebat omnes uendentes et ementes in

templo et mensas nummulariorum et cathedras uendentium

21:13 And said unto them, It is written, My house shall be columbas euertit 21:13 et dicit eis scriptum est domus mea

called the house of prayer; but ye have made it a den of thieves.

domus orationis uocabitur uos autem fecistis eam spelun-

21:14 And the blind and the lame came to him in the tem- cam latronum 21:14 et accesserunt ad eum caeci et claudi

ple; and he healed them.

21:15 And when the chief priests and scribes saw the won- in templo et sanauit eos 21:15 uidentes autem principes sa-

derful things that he did, and the children crying in the temple, and saying, Hosanna to the son of David; they

were sore displeased, cerdotum et scribae mirabilia quae fecit et pueros claman-

tes in templo et dicentes osanna Filio Dauid indignati sunt

21:16 And said unto him, Hearest thou what these say? 21:16 et dixerunt ei audis quid isti dicant Iesus autem dicit

And Jesus saith unto them, Yea; have ye never read, Out of the mouth of babes and sucklings thou hast perfected

praise? eis utique numquam legistis quia ex ore infantium et lac-21:17 And he left them, and went out of the city into tantium perfecisti laudem 21:17 et relictis illis abiit foras

Bethany; and he lodged there.

21:18 Now in the morning as he returned into the city, he extra ciuitatem in Bethaniam ibique mansit 21:18 mane au-

hungered.

21:19 And when he saw a fig tree in the way, he came to it, tem reuertens in ciuitatem esuriit 21:19 et uidens fici arbo-

and found nothing thereon, but leaves only, and said unto it, Let no fruit grow on thee henceforward for ever. And

presently the fig tree withered away. rem unam secus uiam uenit ad eam et nihil inuenit in ea nisi

folia tantum et ait illi numquam ex te fructus nascatur in

21:20 And when the disciples saw it, they marvelled, say- sempiternum et arefacta est continuo ficulnea 21:20 et ui-

ing, How soon is the fig tree withered away!

dentes discipuli mirati sunt dicentes quomodo continuo aruit

21:21 Jesus answered and said unto them, Verily I say unto 21:21 respondens autem Iesus ait eis amen dico uobis si ha-

you, If ye have faith, and doubt not, ye shall not only do this which is done to the fig tree, but also if ye shall say unto this mountain, Be thou removed, and be thou cast into the sea; it shall be done.

bueritis fidem et non haesitaueritis non solum de ficulnea

Euangelium Secundum Matthaeum

facietis sed et si monti huic dixeritis tolle et iacta te in mare fiet 21:22 et omnia quaecumque petieritis in oratione cre- 21:22 And all things, whatsoever ye shall ask in prayer,

believing, ye shall receive.

dentes accipietis 21:23 et cum uenisset in templum acces- 21:23 And when he was come into the temple, the chief

priests and the elders of the people came unto him as he was teaching, and said, By what authority doest thou these

serunt ad eum docentem principes sacerdotum et seniores things? and who gave thee this authority? populi dicentes in qua potestate haec facis et quis tibi dedit hanc potestatem 21:24 respondens Iesus dixit illis inter- 21:24 And Jesus answered and said unto them, I also will

ask you one thing, which if ye tell me, I in like wise will tell you by what authority I do these things.

rogabo uos et ego unum sermonem quem si dixeritis mihi et ego uobis dicam in qua potestate haec facio 21:25 bap- 21:25 The baptism of John, whence was it? from heaven,

or of men? And they reasoned with themselves, saying, If we shall say, From heaven; he will say unto us, Why did

tismum Iohannis unde erat e caelo an ex hominibus at illi ye not then believe him? cogitabant inter se dicentes si dixerimus e caelo dicet nobis quare ergo non credidistis illi 21:26 si autem dixerimus ex 21:26 But if we shall say, Of men; we fear the people; for

all hold John as a prophet.

hominibus timemus turbam omnes enim habent Iohannem sicut prophetam 21:27 et respondentes Iesu dixerunt nesci- 21:27 And they answered Jesus, and said, We cannot tell.

And he said unto them, Neither tell I you by what authority I do these things.

mus ait illis et ipse nec ego dico uobis in qua potestate haec facio 21:28 quid autem uobis uidetur homo habebat duos 21:28 But what think ye? A certain man had two sons; and

he came to the first, and said, Son, go work to day in my vineyard.

filios et accedens ad primum dixit fili uade hodie operare in uinea mea 21:29 ille autem respondens ait nolo postea 21:29 He answered and said, I will not: but afterward he

repented, and went.

autem paenitentia motus abiit 21:30 accedens autem ad al- 21:30 And he came to the second, and said likewise. And

he answered and said, I go, sir: and went not.

terum dixit similiter at ille respondens ait eo domine et non iuit 21:31 quis ex duobus fecit uoluntatem patris dicunt no- 21:31 Whether of them twain did the will of his father?

They say unto him, The first. Jesus saith unto them, Verily I say unto you, That the publicans and the harlots go into

uissimus dicit illis Iesus amen dico uobis quia publicani et the kingdom of God before you. meretrices praecedunt uos in regno Dei 21:32 uenit enim ad 21:32 For John came unto you in the way of righteousness,

and ye believed him not: but the publicans and the harlots believed him: and ye, when ye had seen it, repented not

uos Iohannes in uia iustitiae et non credidistis ei publicani afterward, that ye might believe him. autem et meretrices crediderunt ei uos autem uidentes nec paenitentiam habuistis postea ut crederetis ei 21:33 aliam 21:33 Hear another parable: There was a certain house-

holder, which planted a vineyard, and hedged it round about, and digged a winepress in it, and built a tower, and

parabolam audite homo erat pater familias qui plantauit ui- let it out to husbandmen, and went into a far country: neam et sepem circumdedit ei et fodit in ea torcular et aedificauit turrem et locauit eam agricolis et peregre profectus est 21:34 cum autem tempus fructuum adpropinquasset mi- 21:34 And when the time of the fruit drew near, he sent

his servants to the husbandmen, that they might receive the fruits of it.

sit seruos suos ad agricolas ut acciperent fructus eius 21:35

21:35 And the husbandmen took his servants, and beat one, and killed another, and stoned another.

et agricolae adprehensis seruis eius alium ceciderunt alium occiderunt alium uero lapidauerunt 21:36 iterum misit alios 21:36 Again, he sent other servants more than the first:

and they did unto them likewise.

seruos plures prioribus et fecerunt illis similiter 21:37 no- 21:37 But last of all he sent unto them his son, saying,

They will reverence my son.

Biblia Sacra Vulgata

uissime autem misit ad eos filium suum dicens uerebuntur

21:38 But when the husbandmen saw the son, they said filium meum 21:38 agricolae autem uidentes filium dixe-

among themselves, This is the heir; come, let us kill him, and let us seize on his inheritance.

runt intra se hic est heres uenite occidamus eum et habebi-

21:39 And they caught him, and cast him out of the vine- mus hereditatem eius 21:39 et adprehensum eum eiecerunt

yard, and slew him.

21:40 When the lord therefore of the vineyard cometh, extra uineam et occiderunt 21:40 cum ergo uenerit domi-

what will he do unto those husbandmen?

21:41 They say unto him, He will miserably destroy those nus uineae quid faciet agricolis illis 21:41 aiunt illi malos

wicked men, and will let out his vineyard unto other husbandmen, which shall render him the fruits in their sea-

sons. male perdet et uineam locabit aliis agricolis qui reddant ei 21:42 Jesus saith unto them, Did ye never read in the scrip- fructum temporibus suis 21:42 dicit illis Iesus numquam le-

tures, The stone which the builders rejected, the same is become the head of the corner: this is the Lord’s doing,

and it is marvellous in our eyes? gistis in scripturis lapidem quem reprobauerunt aedificantes

hic factus est in caput anguli a Domino factum est istud et est

21:43 Therefore say I unto you, The kingdom of God shall mirabile in oculis nostris 21:43 ideo dico uobis quia aufere-

be taken from you, and given to a nation bringing forth the fruits thereof.

tur a uobis regnum Dei et dabitur genti facienti fructus eius

21:44 And whosoever shall fall on this stone shall be bro- 21:44 et qui ceciderit super lapidem istum confringetur super

ken: but on whomsoever it shall fall, it will grind him to powder.

quem uero ceciderit conteret eum 21:45 et cum audissent

21:45 And when the chief priests and Pharisees had heard his parables, they perceived that he spake of them.

principes sacerdotum et Pharisaei parabolas eius cognoue-

21:46 But when they sought to lay hands on him, they runt quod de ipsis diceret 21:46 et quaerentes eum tenere

feared the multitude, because they took him for a prophet.

timuerunt turbas quoniam sicut prophetam eum habebant

22:1 And Jesus answered and spake unto them again by 22:1 et respondens Iesus dixit iterum in parabolis eis di-

parables, and said,

22:2 The kingdom of heaven is like unto a certain king, cens 22:2 simile factum est regnum caelorum homini re-

which made a marriage for his son,

22:3 And sent forth his servants to call them that were gi qui fecit nuptias filio suo 22:3 et misit seruos suos uo-

bidden to the wedding: and they would not come.

22:4 Again, he sent forth other servants, saying, Tell them care inuitatos ad nuptias et nolebant uenire 22:4 iterum mi-

which are bidden, Behold, I have prepared my dinner: my oxen and my fatlings are killed, and all things are ready:

come unto the marriage. sit alios seruos dicens dicite inuitatis ecce prandium meum

paraui tauri mei et altilia occisa et omnia parata uenite ad

22:5 But they made light of it, and went their ways, one to nuptias 22:5 illi autem neglexerunt et abierunt alius in uil-

his farm, another to his merchandise:

22:6 And the remnant took his servants, and entreated lam suam alius uero ad negotiationem suam 22:6 reliqui

them spitefully, and slew them.

uero tenuerunt seruos eius et contumelia adfectos occiderunt

22:7 But when the king heard thereof, he was wroth: and 22:7 rex autem cum audisset iratus est et missis exercitibus

he sent forth his armies, and destroyed those murderers, and burned up their city.

suis perdidit homicidas illos et ciuitatem illorum succendit

22:8 Then saith he to his servants, The wedding is ready, 22:8 tunc ait seruis suis nuptiae quidem paratae sunt sed qui

but they which were bidden were not worthy.

22:9 Go ye therefore into the highways, and as many as ye inuitati erant non fuerunt digni 22:9 ite ergo ad exitus uia-

shall find, bid to the marriage.

22:10 So those servants went out into the highways, and rum et quoscumque inueneritis uocate ad nuptias 22:10 et

gathered together all as many as they found, both bad and good: and the wedding was furnished with guests.

egressi serui eius in uias congregauerunt omnes quos inuenerunt malos et bonos et impletae sunt nuptiae discumbentium 22:11 intrauit autem rex ut uideret discumbentes et ui- 22:11 And when the king came in to see the guests, he saw

there a man which had not on a wedding garment:

dit ibi hominem non uestitum ueste nuptiali 22:12 et ait illi 22:12 And he saith unto him, Friend, how camest thou in

hither not having a wedding garment? And he was speechless.

amice quomodo huc intrasti non habens uestem nuptialem at ille obmutuit 22:13 tunc dixit rex ministris ligatis pedi- 22:13 Then said the king to the servants, Bind him hand

and foot, and take him away, and cast him into outer darkness, there shall be weeping and gnashing of teeth.

bus eius et manibus mittite eum in tenebras exteriores ibi erit fletus et stridor dentium 22:14 multi autem sunt uocati 22:14 For many are called, but few are chosen. pauci uero electi 22:15 tunc abeuntes Pharisaei consilium 22:15 Then went the Pharisees, and took counsel how they

might entangle him in his talk.

inierunt ut caperent eum in sermone 22:16 et mittunt ei dis- 22:16 And they sent out unto him their disciples with the

Herodians, saying, Master, we know that thou art true, and teachest the way of God in truth, neither carest thou for

cipulos suos cum Herodianis dicentes magister scimus quia any man: for thou regardest not the person of men. uerax es et uiam Dei in ueritate doces et non est tibi cura

de aliquo non enim respicis personam hominum 22:17 dic 22:17 Tell us therefore, What thinkest thou? Is it lawful to

give tribute unto Caesar, or not?

ergo nobis quid tibi uideatur licet censum dare Caesari an non 22:18 cognita autem Iesus nequitia eorum ait quid me 22:18 But Jesus perceived their wickedness, and said,

Why tempt ye me, ye hypocrites?

temptatis hypocritae 22:19 ostendite mihi nomisma census 22:19 Shew me the tribute money. And they brought unto

him a penny.

at illi obtulerunt ei denarium 22:20 et ait illis Iesus cuius est 22:20 And he saith unto them, Whose is this image and

superscription?

imago haec et suprascriptio 22:21 dicunt ei Caesaris tunc 22:21 They say unto him, Caesar’s. Then saith he unto

them, Render therefore unto Caesar the things which are Caesar’s; and unto God the things that are God’s.

ait illis reddite ergo quae sunt Caesaris Caesari et quae sunt Dei Deo 22:22 et audientes mirati sunt et relicto eo abierunt 22:22 When they had heard these words, they marvelled,

and left him, and went their way.

22:23 in illo die accesserunt ad eum Sadducaei qui dicunt non 22:23 The same day came to him the Sadducees, which

say that there is no resurrection, and asked him,

esse resurrectionem et interrogauerunt eum 22:24 dicentes 22:24 Saying, Master, Moses said, If a man die, having

no children, his brother shall marry his wife, and raise up seed unto his brother.

magister Moses dixit si quis mortuus fuerit non habens filium ut ducat frater eius uxorem illius et suscitet semen fratri suo 22:25 erant autem apud nos septem fratres et primus 22:25 Now there were with us seven brethren: and the

first, when he had married a wife, deceased, and, having no issue, left his wife unto his brother:

uxore ducta defunctus est et non habens semen reliquit uxorem suam fratri suo 22:26 similiter secundus et tertius us- 22:26 Likewise the second also, and the third, unto the

seventh.

que ad septimum 22:27 nouissime autem omnium et mulier 22:27 And last of all the woman died also. defuncta est 22:28 in resurrectione ergo cuius erit de sep- 22:28 Therefore in the resurrection whose wife shall she

be of the seven? for they all had her.

tem uxor omnes enim habuerunt eam 22:29 respondens au- 22:29 Jesus answered and said unto them, Ye do err, not

knowing the scriptures, nor the power of God.

tem Iesus ait illis erratis nescientes scripturas neque uirtutem Dei 22:30 in resurrectione enim neque nubent neque nuben- 22:30 For in the resurrection they neither marry, nor are

given in marriage, but are as the angels of God in heaven.

tur sed sunt sicut angeli Dei in caelo 22:31 de resurrectione 22:31 But as touching the resurrection of the dead, have ye

not read that which was spoken unto you by God, saying,

autem mortuorum non legistis quod dictum est a Deo dicente uobis 22:32 ego sum Deus Abraham et Deus Isaac et 22:32 I am the God of Abraham, and the God of Isaac, and

the God of Jacob? God is not the God of the dead, but of the living.

22:33 And when the multitude heard this, they were as- Deus Iacob non est Deus mortuorum sed uiuentium 22:33 tonished at his doctrine.

22:34 But when the Pharisees had heard that he had put et audientes turbae mirabantur in doctrina eius 22:34 Phari-

the Sadducees to silence, they were gathered together.

saei autem audientes quod silentium inposuisset Sadducaeis

22:35 Then one of them, which was a lawyer, asked him a conuenerunt in unum 22:35 et interrogauit eum unus ex eis

question, tempting him, and saying,

22:36 Master, which is the great commandment in the legis doctor temptans eum 22:36 magister quod est man-

law?

22:37 Jesus said unto him, Thou shalt love the Lord thy datum magnum in lege 22:37 ait illi Iesus diliges Domi-

God with all thy heart, and with all thy soul, and with all thy mind.

num Deum tuum ex toto corde tuo et in tota anima tua et

22:38 This is the first and great commandment. in tota mente tua 22:38 hoc est maximum et primum man-22:39 And the second is like unto it, Thou shalt love thy datum 22:39 secundum autem simile est huic diliges pro-

neighbour as thyself.

22:40 On these two commandments hang all the law and ximum tuum sicut te ipsum 22:40 in his duobus mandatis

the prophets.

22:41 While the Pharisees were gathered together, Jesus uniuersa lex pendet et prophetae 22:41 congregatis autem

asked them,

22:42 Saying, What think ye of Christ? whose son is he? Pharisaeis interrogauit eos Iesus 22:42 dicens quid uobis

They say unto him, The son of David.

22:43 He saith unto them, How then doth David in spirit uidetur de Christo cuius filius est dicunt ei Dauid 22:43 ait

call him Lord, saying,

illis quomodo ergo Dauid in spiritu uocat eum Dominum di-

22:44 The LORD said unto my Lord, Sit thou on my right cens 22:44 dixit Dominus Domino meo sede a dextris meis

hand, till I make thine enemies thy footstool?

22:45 If David then call him Lord, how is he his son? donec ponam inimicos tuos scabillum pedum tuorum 22:45 si ergo Dauid uocat eum Dominum quomodo filius eius est

22:46 And no man was able to answer him a word, nei- 22:46 et nemo poterat respondere ei uerbum neque ausus fuit

ther durst any man from that day forth ask him any more questions.

quisquam ex illa die eum amplius interrogare

23:1 Then spake Jesus to the multitude, and to his disci- 23:1 tunc Iesus locutus est ad turbas et discipulos suos ples,

23:2 Saying The scribes and the Pharisees sit in Moses’ 23:2 dicens super cathedram Mosi sederunt scribae et Phari-

seat:

23:3 All therefore whatsoever they bid you observe, that saei 23:3 omnia ergo quaecumque dixerint uobis seruate et

observe and do; but do not ye after their works: for they say, and do not.

facite secundum opera uero eorum nolite facere dicunt enim

23:4 For they bind heavy burdens and grievous to be et non faciunt 23:4 alligant autem onera grauia et inporta-

borne, and lay them on men’s shoulders; but they themselves will not move them with one of their fingers.

bilia et inponunt in umeros hominum digito autem suo no-

23:5 But all their works they do for to be seen of men: they lunt ea mouere 23:5 omnia uero opera sua faciunt ut uidemake broad their phylacteries, and enlarge the borders of their garments,

antur ab hominibus dilatant enim phylacteria sua et magni-

23:6 And love the uppermost rooms at feasts, and the chief ficant fimbrias 23:6 amant autem primos recubitus in cenis seats in the synagogues,

23:7 And greetings in the markets, and to be called of men, et primas cathedras in synagogis 23:7 et salutationes in foro

Rabbi, Rabbi.

23:8 But be not ye called Rabbi: for one is your Master, et uocari ab hominibus rabbi 23:8 uos autem nolite uocari

even Christ; and all ye are brethren.

rabbi unus enim est magister uester omnes autem uos fra-

23:9 And call no man your father upon the earth: for one tres estis 23:9 et patrem nolite uocare uobis super terram

is your Father, which is in heaven. 23:10 Neither be ye called masters: for one is your Master, unus enim est Pater uester qui in caelis est 23:10 nec uoceeven Christ.

Euangelium Secundum Matthaeum

mini magistri quia magister uester unus est Christus 23:11 23:11 But he that is greatest among you shall be your ser-

vant.

qui maior est uestrum erit minister uester 23:12 qui autem 23:12 And whosoever shall exalt himself shall be abased;

and he that shall humble himself shall be exalted.

se exaltauerit humiliabitur et qui se humiliauerit exaltabitur 23:13 uae autem uobis scribae et Pharisaei hypocritae 23:13 But woe unto you, scribes and Pharisees, hyp-

ocrites! for ye shut up the kingdom of heaven against men: for ye neither go in yourselves, neither suffer ye them that

quia clauditis regnum caelorum ante homines uos enim non are entering to go in. intratis nec introeuntes sinitis intrare 23:14 23:15 uae uo- 23:14 Woe unto you, scribes and Pharisees, hypocrites!

for ye devour widows’ houses, and for a pretence make long prayer: therefore ye shall receive the greater damnation. 23:15 Woe unto you, scribes and Pharisees, hypocrites! for ye compass sea and land to make one proselyte, and when he is made, ye make him twofold more the child of hell than yourselves.

bis scribae et Pharisaei hypocritae quia circuitis mare et aridam ut faciatis unum proselytum et cum fuerit factus facitis

eum filium gehennae duplo quam uos 23:16 uae uobis du- 23:16 Woe unto you, ye blind guides, which say, Whosoever shall swear by the temple, it is nothing; but whosoever shall swear by the gold of the temple, he is a debtor!

ces caeci qui dicitis quicumque iurauerit per templum nihil est qui autem iurauerit in aurum templi debet 23:17 stulti et 23:17 Ye fools and blind: for whether is greater, the gold,

or the temple that sanctifieth the gold?

caeci quid enim maius est aurum an templum quod sanctificat aurum 23:18 et quicumque iurauerit in altari nihil est 23:18 And, Whosoever shall swear by the altar, it is noth-

ing; but whosoever sweareth by the gift that is upon it, he is guilty.

quicumque autem iurauerit in dono quod est super illud de-

bet 23:19 caeci quid enim maius est donum an altare quod 23:19 Ye fools and blind: for whether is greater, the gift,

or the altar that sanctifieth the gift?

sanctificat donum 23:20 qui ergo iurat in altare iurat in eo 23:20 Whoso therefore shall swear by the altar, sweareth

by it, and by all things thereon.

et in omnibus quae super illud sunt 23:21 et qui iurauerit 23:21 And whoso shall swear by the temple, sweareth by

it, and by him that dwelleth therein.

in templo iurat in illo et in eo qui inhabitat in ipso 23:22 et 23:22 And he that shall swear by heaven, sweareth by the

throne of God, and by him that sitteth thereon.

qui iurat in caelo iurat in throno Dei et in eo qui sedet super eum 23:23 uae uobis scribae et Pharisaei hypocritae quia 23:23 Woe unto you, scribes and Pharisees, hypocrites!

for ye pay tithe of mint and anise and cummin, and have omitted the weightier matters of the law, judgment, mercy, and faith: these ought ye to have done, and not to leave the other undone.

decimatis mentam et anethum et cyminum et reliquistis quae grauiora sunt legis iudicium et misericordiam et fidem haec oportuit facere et illa non omittere 23:24 duces caeci exco- 23:24 Ye blind guides, which strain at a gnat, and swallow

a camel.

lantes culicem camelum autem gluttientes 23:25 uae uobis 23:25 Woe unto you, scribes and Pharisees, hypocrites!

for ye make clean the outside of the cup and of the platter, but within they are full of extortion and excess.

scribae et Pharisaei hypocritae quia mundatis quod de foris est calicis et parapsidis intus autem pleni sunt rapina et inmunditia 23:26 Pharisaee caece munda prius quod intus est 23:26 Thou blind Pharisee, cleanse first that which is

within the cup and platter, that the outside of them may be clean also.

calicis et parapsidis ut fiat et id quod de foris est mundum 23:27 uae uobis scribae et Pharisaei hypocritae quia similes 23:27 Woe unto you, scribes and Pharisees, hypocrites!

for ye are like unto whited sepulchres, which indeed appear beautiful outward, but are within full of dead men’s

estis sepulchris dealbatis quae a foris parent hominibus spe- bones, and of all uncleanness. ciosa intus uero plena sunt ossibus mortuorum et omni spurcitia 23:28 sic et uos a foris quidem paretis hominibus iusti 23:28 Even so ye also outwardly appear righteous unto

men, but within ye are full of hypocrisy and iniquity.

intus autem pleni estis hypocrisi et iniquitate 23:29 uae uo- 23:29 Woe unto you, scribes and Pharisees, hypocrites!

because ye build the tombs of the prophets, and garnish the sepulchres of the righteous,

Biblia Sacra Vulgata

bis scribae et Pharisaei hypocritae quia aedificatis sepulchra

23:30 And say, If we had been in the days of our fathers, prophetarum et ornatis monumenta iustorum 23:30 et dici-

we would not have been partakers with them in the blood of the prophets.

tis si fuissemus in diebus patrum nostrorum non essemus so-

23:31 Wherefore ye be witnesses unto yourselves, that ye cii eorum in sanguine prophetarum 23:31 itaque testimonio

are the children of them which killed the prophets.

estis uobismet ipsis quia filii estis eorum qui prophetas oc-

23:32 Fill ye up then the measure of your fathers. ciderunt 23:32 et uos implete mensuram patrum uestrorum 23:33 Ye serpents, ye generation of vipers, how can ye 23:33 serpentes genimina uiperarum quomodo fugietis a iudi-

escape the damnation of hell?

23:34 Wherefore, behold, I send unto you prophets, and cio gehennae 23:34 ideo ecce ego mitto ad uos prophetas et

wise men, and scribes: and some of them ye shall kill and crucify; and some of them shall ye scourge in your syna-

gogues, and persecute them from city to city: sapientes et scribas ex illis occidetis et crucifigetis et ex eis

flagellabitis in synagogis uestris et persequemini de ciuitate

23:35 That upon you may come all the righteous blood in ciuitatem 23:35 ut ueniat super uos omnis sanguis iustus

shed upon the earth, from the blood of righteous Abel unto the blood of Zacharias son of Barachias, whom ye slew

between the temple and the altar. qui effusus est super terram a sanguine Abel iusti usque ad

sanguinem Zacchariae filii Barachiae quem occidistis inter

23:36 Verily I say unto you, All these things shall come templum et altare 23:36 amen dico uobis uenient haec om-

upon this generation.

23:37 O Jerusalem, Jerusalem, thou that killest the nia super generationem istam 23:37 Hierusalem Hierusa-

prophets, and stonest them which are sent unto thee, how often would I have gathered thy children together, even as a hen gathereth her chickens under her wings, and ye would not!

lem quae occidis prophetas et lapidas eos qui ad te missi sunt quotiens uolui congregare filios tuos quemadmodum gallina

23:38 Behold, your house is left unto you desolate. congregat pullos suos sub alas et noluisti 23:38 ecce relin-23:39 For I say unto you, Ye shall not see me henceforth, quitur uobis domus uestra deserta 23:39 dico enim uobis

till ye shall say, Blessed is he that cometh in the name of the Lord.

non me uidebitis amodo donec dicatis benedictus qui uenit in nomine Domini

24:1 And Jesus went out, and departed from the temple: 24:1 et egressus Iesus de templo ibat et accesserunt dis-

and his disciples came to him for to shew him the buildings of the temple.

cipuli eius ut ostenderent ei aedificationes templi 24:2 ipse

24:2 And Jesus said unto them, See ye not all these things? verily I say unto you, There shall not be left here one stone upon another, that shall not be thrown down.

autem respondens dixit eis uidetis haec omnia amen dico uobis non relinquetur hic lapis super lapidem qui non destrua-

24:3 And as he sat upon the mount of Olives, the disciples tur 24:3 sedente autem eo super montem Oliueti accessecame unto him privately, saying, Tell us, when shall these things be? and what shall be the sign of thy coming, and

of the end of the world? runt ad eum discipuli secreto dicentes dic nobis quando haec

erunt et quod signum aduentus tui et consummationis saeculi

24:4 And Jesus answered and said unto them, Take heed 24:4 et respondens Iesus dixit eis uidete ne quis uos seducat

that no man deceive you.

24:5 For many shall come in my name, saying, I am 24:5 multi enim uenient in nomine meo dicentes ego sum

Christ; and shall deceive many.

24:6 And ye shall hear of wars and rumours of wars: see Christus et multos seducent 24:6 audituri autem estis proe-

that ye be not troubled: for all these things must come to pass, but the end is not yet.

lia et opiniones proeliorum uidete ne turbemini oportet enim

24:7 For nation shall rise against nation, and kingdom haec fieri sed nondum est finis 24:7 consurget enim gens in

against kingdom: and there shall be famines, and pestilences, and earthquakes, in divers places.

Euangelium Secundum Matthaeum

gentem et regnum in regnum et erunt pestilentiae et fames et terraemotus per loca 24:8 haec autem omnia initia sunt do- 24:8 All these are the beginning of sorrows. lorum 24:9 tunc tradent uos in tribulationem et occident uos 24:9 Then shall they deliver you up to be afflicted, and

shall kill you: and ye shall be hated of all nations for my name’s sake.

et eritis odio omnibus gentibus propter nomen meum 24:10

24:10 And then shall many be offended, and shall betray one another, and shall hate one another.

et tunc scandalizabuntur multi et inuicem tradent et odio habebunt inuicem 24:11 et multi pseudoprophetae surgent et 24:11 And many false prophets shall rise, and shall de-

ceive many.

seducent multos 24:12 et quoniam abundabit iniquitas refri- 24:12 And because iniquity shall abound, the love of many

shall wax cold.

gescet caritas multorum 24:13 qui autem permanserit usque 24:13 But he that shall endure unto the end, the same shall

be saved.

in finem hic saluus erit 24:14 et praedicabitur hoc euange- 24:14 And this gospel of the kingdom shall be preached in

all the world for a witness unto all nations; and then shall the end come.

lium regni in uniuerso orbe in testimonium omnibus genti-

bus et tunc ueniet consummatio 24:15 cum ergo uideritis 24:15 When ye therefore shall see the abomination of desolation, spoken of by Daniel the prophet, stand in the holy place, (whoso readeth, let him understand:)

abominationem desolationis quae dicta est a Danihelo propheta stantem in loco sancto qui legit intellegat 24:16 tunc 24:16 Then let them which be in Judaea flee into the

mountains:

qui in Iudaea sunt fugiant ad montes 24:17 et qui in tecto 24:17 Let him which is on the housetop not come down to

take any thing out of his house:

non descendat tollere aliquid de domo sua 24:18 et qui in 24:18 Neither let him which is in the field return back to

take his clothes.

agro non reuertatur tollere tunicam suam 24:19 uae autem 24:19 And woe unto them that are with child, and to them

that give suck in those days!

praegnatibus et nutrientibus in illis diebus 24:20 orate au- 24:20 But pray ye that your flight be not in the winter,

neither on the sabbath day:

tem ut non fiat fuga uestra hieme uel sabbato 24:21 erit 24:21 For then shall be great tribulation, such as was not since the beginning of the world to this time, no, nor ever shall be.

enim tunc tribulatio magna qualis non fuit ab initio mundi usque modo neque fiet 24:22 et nisi breuiati fuissent dies 24:22 And except those days should be shortened, there

should no flesh be saved: but for the elect’s sake those days shall be shortened.

illi non fieret salua omnis caro sed propter electos breuiabuntur dies illi 24:23 tunc si quis uobis dixerit ecce hic Ch- 24:23 Then if any man shall say unto you, Lo, here is

Christ, or there; believe it not.

ristus aut illic nolite credere 24:24 surgent enim pseudo- 24:24 For there shall arise false Christs, and false

prophets, and shall shew great signs and wonders; insomuch that, if it were possible, they shall deceive the very

christi et pseudoprophetae et dabunt signa magna et prod- elect. iiia ita ut in errorem inducantur si fieri potest etiam electi 24:25 ecce praedixi uobis 24:26 si ergo dixerint uobis ecce 24:25 Behold, I have told you before.

24:26 Wherefore if they shall say unto you, Behold, he is in the desert; go not forth: behold, he is in the secret chambers; believe it not.

in deserto est nolite exire ecce in penetrabilibus nolite credere 24:27 sicut enim fulgur exit ab oriente et paret usque 24:27 For as the lightning cometh out of the east, and

shineth even unto the west; so shall also the coming of the Son of man be.

in occidente ita erit et aduentus Filii hominis 24:28 ubicum-

24:28 For wheresoever the carcase is, there will the eagles be gathered together.

que fuerit corpus illuc congregabuntur aquilae 24:29 statim 24:29 Immediately after the tribulation of those days shall

the sun be darkened, and the moon shall not give her light, and the stars shall fall from heaven, and the powers of the

autem post tribulationem dierum illorum sol obscurabitur et heavens shall be shaken: luna non dabit lumen suum et stellae cadent de caelo et uirtutes caelorum commouebuntur 24:30 et tunc parebit signum 24:30 And then shall appear the sign of the Son of man

in heaven: and then shall all the tribes of the earth mourn, and they shall see the Son of man coming in the clouds of heaven with power and great glory.

Filii hominis in caelo et tunc plangent omnes tribus terrae et uidebunt Filium hominis uenientem in nubibus caeli cum

24:31 And he shall send his angels with a great sound of a uirtute multa et maiestate 24:31 et mittet angelos suos cum

trumpet, and they shall gather together his elect from the four winds, from one end of heaven to the other.

tuba et uoce magna et congregabunt electos eius a quattuor

24:32 Now learn a parable of the fig tree; When his branch uentis a summis caelorum usque ad terminos eorum 24:32 is yet tender, and putteth forth leaves, ye know that summer is nigh:

ab arbore autem fici discite parabolam cum iam ramus eius

24:33 So likewise ye, when ye shall see all these things, tener fuerit et folia nata scitis quia prope est aestas 24:33

know that it is near, even at the doors.

ita et uos cum uideritis haec omnia scitote quia prope est in

24:34 Verily I say unto you, This generation shall not pass, ianuis 24:34 amen dico uobis quia non praeteribit haec getill all these things be fulfilled.

24:35 Heaven and earth shall pass away, but my words neratio donec omnia haec fiant 24:35 caelum et terra trans-

shall not pass away.

24:36 But of that day and hour knoweth no man, no, not ibunt uerba uero mea non praeteribunt 24:36 de die autem

the angels of heaven, but my Father only.

illa et hora nemo scit neque angeli caelorum nisi Pater so-

24:37 But as the days of Noe were, so shall also the com- lus 24:37 sicut autem in diebus Noe ita erit et aduentus Filii

ing of the Son of man be.

24:38 For as in the days that were before the flood they hominis 24:38 sicut enim erant in diebus ante diluuium co-

were eating and drinking, marrying and giving in marriage, until the day that Noe entered into the ark,

medentes et bibentes nubentes et nuptum tradentes usque ad

24:39 And knew not until the flood came, and took them eum diem quo introiuit in arcam Noe 24:39 et non cognoue-

all away; so shall also the coming of the Son of man be.

runt donec uenit diluuium et tulit omnes ita erit et aduentus

24:40 Then shall two be in the field; the one shall be taken, Filii hominis 24:40 tunc duo erunt in agro unus adsumetur and the other left.

24:41 Two women shall be grinding at the mill; the one et unus relinquetur 24:41 duae molentes in mola una ad-

shall be taken, and the other left.

24:42 Watch therefore: for ye know not what hour your sumetur et una relinquetur 24:42 uigilate ergo quia nesci-

Lord doth come.

24:43 But know this, that if the goodman of the house had tis qua hora Dominus uester uenturus sit 24:43 illud autem

known in what watch the thief would come, he would have watched, and would not have suffered his house to be bro-

ken up. scitote quoniam si sciret pater familias qua hora fur uenturus esset uigilaret utique et non sineret perfodiri domum

24:44 Therefore be ye also ready: for in such an hour as suam 24:44 ideoque et uos estote parati quia qua nescitis

ye think not the Son of man cometh.

24:45 Who then is a faithful and wise servant, whom his hora Filius hominis uenturus est 24:45 quis putas est fidelis

lord hath made ruler over his household, to give them meat in due season?

seruus et prudens quem constituit dominus suus supra fami-

24:46 Blessed is that servant, whom his lord when he liam suam ut det illis cibum in tempore 24:46 beatus ille

cometh shall find so doing.

seruus quem cum uenerit dominus eius inuenerit sic facien-

24:47 Verily I say unto you, That he shall make him ruler tem 24:47 amen dico uobis quoniam super omnia bona sua over all his goods. 24:48 But and if that evil servant shall say in his heart, My constituet eum 24:48 si autem dixerit malus seruus ille in lord delayeth his coming; 24:49 And shall begin to smite his fellowservants, and to corde suo moram facit dominus meus uenire 24:49 et coeeat and drink with the drunken;

perit percutere conseruos suos manducet autem et bibat cum

24:50 The lord of that servant shall come in a day when he ebriis 24:50 ueniet dominus serui illius in die qua non spe-

looketh not for him, and in an hour that he is not aware of,

Euangelium Secundum Matthaeum

rat et hora qua ignorat 24:51 et diuidet eum partemque eius 24:51 And shall cut him asunder, and appoint him his por-

tion with the hypocrites: there shall be weeping and gnashing of teeth.

ponet cum hypocritis illic erit fletus et stridor dentium

25:1 tunc simile erit regnum caelorum decem uirginibus 25:1 Then shall the kingdom of heaven be likened unto ten

virgins, which took their lamps, and went forth to meet the bridegroom.

quae accipientes lampadas suas exierunt obuiam sponso et sponsae 25:2 quinque autem ex eis erant fatuae et quin- 25:2 And five of them were wise, and five were foolish. que prudentes 25:3 sed quinque fatuae acceptis lampadi- 25:3 They that were foolish took their lamps, and took no

oil with them:

bus non sumpserunt oleum secum 25:4 prudentes uero ac- 25:4 But the wise took oil in their vessels with their lamps. ceperunt oleum in uasis suis cum lampadibus 25:5 moram 25:5 While the bridegroom tarried, they all slumbered and

slept.

autem faciente sponso dormitauerunt omnes et dormierunt 25:6 media autem nocte clamor factus est ecce sponsus ue- 25:6 And at midnight there was a cry made, Behold, the

bridegroom cometh; go ye out to meet him.

nit exite obuiam ei 25:7 tunc surrexerunt omnes uirgines il- 25:7 Then all those virgins arose, and trimmed their lamps. lae et ornauerunt lampades suas 25:8 fatuae autem sapien- 25:8 And the foolish said unto the wise, Give us of your

oil; for our lamps are gone out.

tibus dixerunt date nobis de oleo uestro quia lampades nostrae extinguntur 25:9 responderunt prudentes dicentes ne 25:9 But the wise answered, saying, Not so; lest there be

not enough for us and you: but go ye rather to them that sell, and buy for yourselves.

forte non sufficiat nobis et uobis ite potius ad uendentes et emite uobis 25:10 dum autem irent emere uenit sponsus et 25:10 And while they went to buy, the bridegroom came;

and they that were ready went in with him to the marriage: and the door was shut.

quae paratae erant intrauerunt cum eo ad nuptias et clausa est ianua 25:11 nouissime ueniunt et reliquae uirgines di- 25:11 Afterward came also the other virgins, saying, Lord,

Lord, open to us.

centes domine domine aperi nobis 25:12 at ille respondens 25:12 But he answered and said, Verily I say unto you, I

know you not.

ait amen dico uobis nescio uos 25:13 uigilate itaque quia 25:13 Watch therefore, for ye know neither the day nor the

hour wherein the Son of man cometh.

nescitis diem neque horam 25:14 sicut enim homo profici- 25:14 For the kingdom of heaven is as a man travelling

into a far country, who called his own servants, and delivered unto them his goods.

scens uocauit seruos suos et tradidit illis bona sua 25:15 et

25:15 And unto one he gave five talents, to another two, and to another one; to every man according to his several ability; and straightway took his journey.

uni dedit quinque talenta alii autem duo alii uero unum unicuique secundum propriam uirtutem et profectus est statim 25:16 abiit autem qui quinque talenta acceperat et operatus 25:16 Then he that had received the five talents went and

traded with the same, and made them other five talents.

est in eis et lucratus est alia quinque 25:17 similiter qui duo 25:17 And likewise he that had received two, he also

gained other two.

acceperat lucratus est alia duo 25:18 qui autem unum ac- 25:18 But he that had received one went and digged in the

earth, and hid his lord’s money.

ceperat abiens fodit in terra et abscondit pecuniam domini sui 25:19 post multum uero temporis uenit dominus seruo- 25:19 After a long time the lord of those servants cometh,

and reckoneth with them.

rum illorum et posuit rationem cum eis 25:20 et accedens 25:20 And so he that had received five talents came and

brought other five talents, saying, Lord, thou deliveredst unto me five talents: behold, I have gained beside them

qui quinque talenta acceperat obtulit alia quinque talenta di- five talents more. cens domine quinque talenta mihi tradidisti ecce alia quinque superlucratus sum 25:21 ait illi dominus eius euge bone 25:21 His lord said unto him, Well done, thou good and

faithful servant: thou hast been faithful over a few things, I will make thee ruler over many things: enter thou into the joy of thy lord.

Biblia Sacra Vulgata

serue et fidelis quia super pauca fuisti fidelis super multa te

25:22 He also that had received two talents came and said, constituam intra in gaudium domini tui 25:22 accessit au-

Lord, thou deliveredst unto me two talents: behold, I have gained two other talents beside them.

tem et qui duo talenta acceperat et ait domine duo talenta

25:23 His lord said unto him, Well done, good and faithful tradidisti mihi ecce alia duo lucratus sum 25:23 ait illi do-

servant; thou hast been faithful over a few things, I will make thee ruler over many things: enter thou into the joy

of thy lord. minus eius euge serue bone et fidelis quia super pauca fuisti

fidelis supra multa te constituam intra in gaudium domini tui

25:24 Then he which had received the one talent came and 25:24 accedens autem et qui unum talentum acceperat ait do-

said, Lord, I knew thee that thou art an hard man, reaping where thou hast not sown, and gathering where thou hast

not strawed: mine scio quia homo durus es metis ubi non seminasti et 25:25 And I was afraid, and went and hid thy talent in the congregas ubi non sparsisti 25:25 et timens abii et abscondi

earth: lo, there thou hast that is thine.

25:26 His lord answered and said unto him, Thou wicked talentum tuum in terra ecce habes quod tuum est 25:26 re-

and slothful servant, thou knewest that I reap where I sowed not, and gather where I have not strawed:

spondens autem dominus eius dixit ei serue male et piger sciebas quia meto ubi non semino et congrego ubi non sparsi

25:27 Thou oughtest therefore to have put my money to 25:27 oportuit ergo te mittere pecuniam meam nummulariis

the exchangers, and then at my coming I should have received mine own with usury.

et ueniens ego recepissem utique quod meum est cum usura

25:28 Take therefore the talent from him, and give it unto 25:28 tollite itaque ab eo talentum et date ei qui habet de-

him which hath ten talents.

25:29 For unto every one that hath shall be given, and he cem talenta 25:29 omni enim habenti dabitur et abundabit

shall have abundance: but from him that hath not shall be taken away even that which he hath.

ei autem qui non habet et quod uidetur habere auferetur ab

25:30 And cast ye the unprofitable servant into outer dark- eo 25:30 et inutilem seruum eicite in tenebras exteriores il-

ness: there shall be weeping and gnashing of teeth.

25:31 When the Son of man shall come in his glory, and all lic erit fletus et stridor dentium 25:31 cum autem uenerit

the holy angels with him, then shall he sit upon the throne of his glory:

Filius hominis in maiestate sua et omnes angeli cum eo tunc

25:32 And before him shall be gathered all nations: and sedebit super sedem maiestatis suae 25:32 et congregabun-

he shall separate them one from another, as a shepherd divideth his sheep from the goats:

tur ante eum omnes gentes et separabit eos ab inuicem sicut

25:33 And he shall set the sheep on his right hand, but the pastor segregat oues ab hedis 25:33 et statuet oues quidem a

goats on the left.

25:34 Then shall the King say unto them on his right hand, dextris suis hedos autem a sinistris 25:34 tunc dicet rex his

Come, ye blessed of my Father, inherit the kingdom prepared for you from the foundation of the world:

qui a dextris eius erunt uenite benedicti Patris mei possidete

25:35 For I was an hungred, and ye gave me meat: I was paratum uobis regnum a constitutione mundi 25:35 esuriui

thirsty, and ye gave me drink: I was a stranger, and ye took me in:

enim et dedistis mihi manducare sitiui et dedistis mihi bibere

25:36 Naked, and ye clothed me: I was sick, and ye visited hospes eram et collexistis me 25:36 nudus et operuistis me

me: I was in prison, and ye came unto me.

infirmus et uisitastis me in carcere eram et uenistis ad me

25:37 Then shall the righteous answer him, saying, Lord, 25:37 tunc respondebunt ei iusti dicentes Domine quando te

when saw we thee an hungred, and fed thee? or thirsty, and gave thee drink?

uidimus esurientem et pauimus sitientem et dedimus tibi po-

25:38 When saw we thee a stranger, and took thee in? or tum 25:38 quando autem te uidimus hospitem et colleximus

naked, and clothed thee?

25:39 Or when saw we thee sick, or in prison, and came te aut nudum et cooperuimus 25:39 aut quando te uidimus

unto thee?

Euangelium Secundum Matthaeum

infirmum aut in carcere et uenimus ad te 25:40 et respon- 25:40 And the King shall answer and say unto them, Verily I say unto you, Inasmuch as ye have done it unto one of the least of these my brethren, ye have done it unto me.

dens rex dicet illis amen dico uobis quamdiu fecistis uni de his fratribus meis minimis mihi fecistis 25:41 tunc dicet et 25:41 Then shall he say also unto them on the left hand,

Depart from me, ye cursed, into everlasting fire, prepared for the devil and his angels:

his qui a sinistris erunt discedite a me maledicti in ignem aeternum qui paratus est diabolo et angelis eius 25:42 esu- 25:42 For I was an hungred, and ye gave me no meat: I

was thirsty, and ye gave me no drink:

riui enim et non dedistis mihi manducare sitiui et non dedistis mihi potum 25:43 hospes eram et non collexistis me 25:43 I was a stranger, and ye took me not in: naked, and

ye clothed me not: sick, and in prison, and ye visited me not.

nudus et non operuistis me infirmus et in carcere et non uisitastis me 25:44 tunc respondebunt et ipsi dicentes Domine 25:44 Then shall they also answer him, saying, Lord,

when saw we thee an hungred, or athirst, or a stranger, or naked, or sick, or in prison, and did not minister unto

quando te uidimus esurientem aut sitientem aut hospitem aut thee? nudum aut infirmum uel in carcere et non ministrauimus tibi 25:45 tunc respondebit illis dicens amen dico uobis quamdiu 25:45 Then shall he answer them, saying, Verily I say unto

you, Inasmuch as ye did it not to one of the least of these, ye did it not to me.

non fecistis uni de minoribus his nec mihi fecistis 25:46 et

25:46 And these shall go away into everlasting punishment: but the righteous into life eternal.

ibunt hii in supplicium aeternum iusti autem in uitam aeternam

26:1 et factum est cum consummasset Iesus sermones hos 26:1 And it came to pass, when Jesus had finished all these

sayings, he said unto his disciples,

omnes dixit discipulis suis 26:2 scitis quia post biduum 26:2 Ye know that after two days is the feast of the

passover, and the Son of man is betrayed to be crucified.

pascha fiet et Filius hominis tradetur ut crucifigatur 26:3 26:3 Then assembled together the chief priests, and the

scribes, and the elders of the people, unto the palace of the high priest, who was called Caiaphas,

tunc congregati sunt principes sacerdotum et seniores populi in atrium principis sacerdotum qui dicebatur Caiaphas 26:4 et consilium fecerunt ut Iesum dolo tenerent et occide- 26:4 And consulted that they might take Jesus by subtilty,

and kill him.

rent 26:5 dicebant autem non in die festo ne forte tumultus 26:5 But they said, Not on the feast day, lest there be an

uproar among the people.

fieret in populo 26:6 cum autem esset Iesus in Bethania in 26:6 Now when Jesus was in Bethany, in the house of Si-

mon the leper,

domo Simonis leprosi 26:7 accessit ad eum mulier habens 26:7 There came unto him a woman having an alabaster

box of very precious ointment, and poured it on his head, as he sat at meat.

alabastrum unguenti pretiosi et effudit super caput ipsius recumbentis 26:8 uidentes autem discipuli indignati sunt di- 26:8 But when his disciples saw it, they had indignation,

saying, To what purpose is this waste?

centes ut quid perditio haec 26:9 potuit enim istud uenun- 26:9 For this ointment might have been sold for much, and

given to the poor.

dari multo et dari pauperibus 26:10 sciens autem Iesus ait 26:10 When Jesus understood it, he said unto them, Why

trouble ye the woman? for she hath wrought a good work upon me.

illis quid molesti estis mulieri opus bonum operata est in me 26:11 nam semper pauperes habetis uobiscum me au- 26:11 For ye have the poor always with you; but me ye

have not always.

tem non semper habetis 26:12 mittens enim haec unguen- 26:12 For in that she hath poured this ointment on my

body, she did it for my burial.

tum hoc in corpus meum ad sepeliendum me fecit 26:13 26:13 Verily I say unto you, Wheresoever this gospel shall be preached in the whole world, there shall also this, that this woman hath done, be told for a memorial of her.

amen dico uobis ubicumque praedicatum fuerit hoc euange-

Biblia Sacra Vulgata

lium in toto mundo dicetur et quod haec fecit in memoriam

26:14 Then one of the twelve, called Judas Iscariot, went eius 26:14 tunc abiit unus de duodecim qui dicitur Iudas unto the chief priests,

26:15 And said unto them, What will ye give me, and I Scarioth ad principes sacerdotum 26:15 et ait illis quid uul-

will deliver him unto you? And they covenanted with him for thirty pieces of silver.

tis mihi dare et ego uobis eum tradam at illi constituerunt ei

26:16 And from that time he sought opportunity to betray triginta argenteos 26:16 et exinde quaerebat oportunitatem

him.

26:17 Now the first day of the feast of unleavened bread ut eum traderet 26:17 prima autem azymorum accesserunt

the disciples came to Jesus, saying unto him, Where wilt thou that we prepare for thee to eat the passover?

discipuli ad Iesum dicentes ubi uis paremus tibi comedere

26:18 And he said, Go into the city to such a man, and say pascha 26:18 at Iesus dixit ite in ciuitatem ad quendam et

unto him, The Master saith, My time is at hand; I will keep the passover at thy house with my disciples.

dicite ei magister dicit tempus meum prope est apud te facio

26:19 And the disciples did as Jesus had appointed them; pascha cum discipulis meis 26:19 et fecerunt discipuli sic-

and they made ready the passover.

26:20 Now when the even was come, he sat down with the ut constituit illis Iesus et parauerunt pascha 26:20 uespere

twelve.

26:21 And as they did eat, he said, Verily I say unto you, autem facto discumbebat cum duodecim discipulis 26:21 et

that one of you shall betray me.

edentibus illis dixit amen dico uobis quia unus uestrum me

26:22 And they were exceeding sorrowful, and began ev- traditurus est 26:22 et contristati ualde coeperunt singuli di-

ery one of them to say unto him, Lord, is it I? 26:23 And he answered and said, He that dippeth his hand cere numquid ego sum Domine 26:23 at ipse respondens with me in the dish, the same shall betray me.

ait qui intinguit mecum manum in parapside hic me tradet

26:24 The Son of man goeth as it is written of him: but 26:24 Filius quidem hominis uadit sicut scriptum est de illo

woe unto that man by whom the Son of man is betrayed! it had been good for that man if he had not been born.

uae autem homini illi per quem Filius hominis traditur bo-

26:25 Then Judas, which betrayed him, answered and said, num erat ei si natus non fuisset homo ille 26:25 respondens

Master, is it I? He said unto him, Thou hast said.

autem Iudas qui tradidit eum dixit numquid ego sum rabbi

26:26 And as they were eating, Jesus took bread, and ait illi tu dixisti 26:26 cenantibus autem eis accepit Iesus

blessed it, and brake it, and gave it to the disciples, and said, Take, eat; this is my body.

panem et benedixit ac fregit deditque discipulis suis et ait ac-

26:27 And he took the cup, and gave thanks, and gave it to cipite et comedite hoc est corpus meum 26:27 et accipiens

them, saying, Drink ye all of it;

calicem gratias egit et dedit illis dicens bibite ex hoc omnes

26:28 For this is my blood of the new testament, which is 26:28 hic est enim sanguis meus noui testamenti qui pro mul-

shed for many for the remission of sins.

26:29 But I say unto you, I will not drink henceforth of tis effunditur in remissionem peccatorum 26:29 dico autem

this fruit of the vine, until that day when I drink it new with you in my Father’s kingdom.

uobis non bibam amodo de hoc genimine uitis usque in diem illum cum illud bibam uobiscum nouum in regno Patris mei

26:30 And when they had sung an hymn, they went out 26:30 et hymno dicto exierunt in montem Oliueti 26:31 tunc

into the mount of Olives. 26:31 Then saith Jesus unto them, All ye shall be offended because of me this night: for it is written, I will smite the shepherd, and the sheep of the flock shall be scattered abroad.

dicit illis Iesus omnes uos scandalum patiemini in me in ista nocte scriptum est enim percutiam pastorem et dispergentur

26:32 But after I am risen again, I will go before you into oues gregis 26:32 postquam autem resurrexero praecedam

Galilee.

26:33 Peter answered and said unto him, Though all men uos in Galilaeam 26:33 respondens autem Petrus ait illi et si

shall be offended because of thee, yet will I never be offended.

Euangelium Secundum Matthaeum

omnes scandalizati fuerint in te ego numquam scandalizabor 26:34 ait illi Iesus amen dico tibi quia in hac nocte antequam 26:34 Jesus said unto him, Verily I say unto thee, That this

night, before the cock crow, thou shalt deny me thrice.

gallus cantet ter me negabis 26:35 ait illi Petrus etiam si 26:35 Peter said unto him, Though I should die with thee,

yet will I not deny thee. Likewise also said all the disciples.

oportuerit me mori tecum non te negabo similiter et omnes discipuli dixerunt 26:36 tunc uenit Iesus cum illis in uillam 26:36 Then cometh Jesus with them unto a place called

Gethsemane, and saith unto the disciples, Sit ye here, while I go and pray yonder.

quae dicitur Gethsemani et dixit discipulis suis sedete hic donec uadam illuc et orem 26:37 et adsumpto Petro et duo- 26:37 And he took with him Peter and the two sons of

Zebedee, and began to be sorrowful and very heavy.

bus filiis Zebedaei coepit contristari et maestus esse 26:38 26:38 Then saith he unto them, My soul is exceeding sor-

rowful, even unto death: tarry ye here, and watch with me.

tunc ait illis tristis est anima mea usque ad mortem sustinete hic et uigilate mecum 26:39 et progressus pusillum proci- 26:39 And he went a little farther, and fell on his face, and

prayed, saying, O my Father, if it be possible, let this cup pass from me: nevertheless not as I will, but as thou wilt.

dit in faciem suam orans et dicens mi Pater si possibile est transeat a me calix iste uerumtamen non sicut ego uolo sed sicut tu 26:40 et uenit ad discipulos et inuenit eos dormien- 26:40 And he cometh unto the disciples, and findeth them

asleep, and saith unto Peter, What, could ye not watch with me one hour?

tes et dicit Petro sic non potuistis una hora uigilare mecum 26:41 uigilate et orate ut non intretis in temptationem spiri- 26:41 Watch and pray, that ye enter not into temptation:

the spirit indeed is willing, but the flesh is weak.

tus quidem promptus est caro autem infirma 26:42 iterum 26:42 He went away again the second time, and prayed,

saying, O my Father, if this cup may not pass away from me, except I drink it, thy will be done.

secundo abiit et orauit dicens Pater mi si non potest hic calix transire nisi bibam illum fiat uoluntas tua 26:43 et uenit 26:43 And he came and found them asleep again: for their

eyes were heavy.

iterum et inuenit eos dormientes erant enim oculi eorum grauati 26:44 et relictis illis iterum abiit et orauit tertio eundem 26:44 And he left them, and went away again, and prayed

the third time, saying the same words.

sermonem dicens 26:45 tunc uenit ad discipulos suos et di- 26:45 Then cometh he to his disciples, and saith unto

them, Sleep on now, and take your rest: behold, the hour is at hand, and the Son of man is betrayed into the hands

cit illis dormite iam et requiescite ecce adpropinquauit hora of sinners. et Filius hominis traditur in manus peccatorum 26:46 sur- 26:46 Rise, let us be going: behold, he is at hand that doth

betray me.

gite eamus ecce adpropinquauit qui me tradit 26:47 adhuc 26:47 And while he yet spake, lo, Judas, one of the twelve,

came, and with him a great multitude with swords and staves, from the chief priests and elders of the people.

ipso loquente ecce Iudas unus de duodecim uenit et cum eo turba multa cum gladiis et fustibus a principibus sacerdotum et senioribus populi 26:48 qui autem tradidit eum dedit illis 26:48 Now he that betrayed him gave them a sign, saying,

Whomsoever I shall kiss, that same is he: hold him fast.

signum dicens quemcumque osculatus fuero ipse est tenete eum 26:49 et confestim accedens ad Iesum dixit haue rabbi 26:49 And forthwith he came to Jesus, and said, Hail, mas-

ter; and kissed him.

et osculatus est eum 26:50 dixitque illi Iesus amice ad quod 26:50 And Jesus said unto him, Friend, wherefore art thou

come? Then came they, and laid hands on Jesus and took him.

uenisti tunc accesserunt et manus iniecerunt in Iesum et tenuerunt eum 26:51 et ecce unus ex his qui erant cum Iesu 26:51 And, behold, one of them which were with Jesus

stretched out his hand, and drew his sword, and struck a servant of the high priest’s, and smote off his ear.

extendens manum exemit gladium suum et percutiens ser-

Biblia Sacra Vulgata

26:52 Then said Jesus unto him, Put up again thy sword uum principis sacerdotum amputauit auriculam eius 26:52

into his place: for all they that take the sword shall perish with the sword.

tunc ait illi Iesus conuerte gladium tuum in locum suum om-

26:53 Thinkest thou that I cannot now pray to my Father, nes enim qui acceperint gladium gladio peribunt 26:53 an

and he shall presently give me more than twelve legions of angels?

putas quia non possum rogare Patrem meum et exhibebit

26:54 But how then shall the scriptures be fulfilled, that mihi modo plus quam duodecim legiones angelorum 26:54 thus it must be?

quomodo ergo implebuntur scripturae quia sic oportet fieri

26:55 In that same hour said Jesus to the multitudes, Are 26:55 in illa hora dixit Iesus turbis tamquam ad latronem ex-

ye come out as against a thief with swords and staves for to take me? I sat daily with you teaching in the temple,

and ye laid no hold on me. istis cum gladiis et fustibus conprehendere me cotidie apud

26:56 But all this was done, that the scriptures of the uos sedebam docens in templo et non me tenuistis 26:56 prophets might be fulfilled. Then all the disciples forsook him, and fled.

hoc autem totum factum est ut implerentur scripturae pro-

26:57 And they that had laid hold on Jesus led him away to phetarum tunc discipuli omnes relicto eo fugerunt 26:57 at

Caiaphas the high priest, where the scribes and the elders were assembled.

illi tenentes Iesum duxerunt ad Caiaphan principem sacerdo-

26:58 But Peter followed him afar off unto the high priest’s tum ubi scribae et seniores conuenerant 26:58 Petrus autem

palace, and went in, and sat with the servants, to see the end.

sequebatur eum a longe usque in atrium principis sacerdotum et ingressus intro sedebat cum ministris ut uideret fi-

26:59 Now the chief priests, and elders, and all the coun- nem 26:59 principes autem sacerdotum et omne concilium cil, sought false witness against Jesus, to put him to death;

quaerebant falsum testimonium contra Iesum ut eum morti

26:60 But found none: yea, though many false witnesses traderent 26:60 et non inuenerunt cum multi falsi testes ac-

came, yet found they none. At the last came two false witnesses,

cessissent nouissime autem uenerunt duo falsi testes 26:61

26:61 And said, This fellow said, I am able to destroy the temple of God, and to build it in three days.

et dixerunt hic dixit possum destruere templum Dei et post

26:62 And the high priest arose, and said unto him, An- triduum aedificare illud 26:62 et surgens princeps sacerdo-

swerest thou nothing? what is it which these witness against thee?

tum ait illi nihil respondes ad ea quae isti aduersum te tes-

26:63 But Jesus held his peace, And the high priest an- tificantur 26:63 Iesus autem tacebat et princeps sacerdotum

swered and said unto him, I adjure thee by the living God, that thou tell us whether thou be the Christ, the Son of

God. ait illi adiuro te per Deum uiuum ut dicas nobis si tu es Ch-26:64 Jesus saith unto him, Thou hast said: nevertheless I ristus Filius Dei 26:64 dicit illi Iesus tu dixisti uerumtamen

say unto you, Hereafter shall ye see the Son of man sitting on the right hand of power, and coming in the clouds of

heaven. dico uobis amodo uidebitis Filium hominis sedentem a de-26:65 Then the high priest rent his clothes, saying, He xtris uirtutis et uenientem in nubibus caeli 26:65 tunc prin-

hath spoken blasphemy; what further need have we of witnesses? behold, now ye have heard his blasphemy.

ceps sacerdotum scidit uestimenta sua dicens blasphemauit quid adhuc egemus testibus ecce nunc audistis blasphemiam

26:66 What think ye? They answered and said, He is 26:66 quid uobis uidetur at illi respondentes dixerunt reus est

guilty of death.

26:67 Then did they spit in his face, and buffeted him; and mortis 26:67 tunc expuerunt in faciem eius et colaphis eum

others smote him with the palms of their hands,

26:68 Saying, Prophesy unto us, thou Christ, Who is he ceciderunt alii autem palmas in faciem ei dederunt 26:68

that smote thee?

dicentes prophetiza nobis Christe quis est qui te percussit

Euangelium Secundum Matthaeum

26:69 Petrus uero sedebat foris in atrio et accessit ad eum 26:69 Now Peter sat without in the palace: and a damsel

came unto him, saying, Thou also wast with Jesus of Galilee.

una ancilla dicens et tu cum Iesu Galilaeo eras 26:70 at ille

26:70 But he denied before them all, saying, I know not what thou sayest.

negauit coram omnibus dicens nescio quid dicis 26:71 ex- 26:71 And when he was gone out into the porch, another

maid saw him, and said unto them that were there, This fellow was also with Jesus of Nazareth.

eunte autem illo ianuam uidit eum alia et ait his qui erant ibi et hic erat cum Iesu Nazareno

26:72 et iterum negauit cum iuramento quia non noui homi- 26:72 And again he denied with an oath, I do not know the

man.

nem 26:73 et post pusillum accesserunt qui stabant et dixe- 26:73 And after a while came unto him they that stood by,

and said to Peter, Surely thou also art one of them; for thy speech bewrayeth thee.

runt Petro uere et tu ex illis es nam et loquella tua manifestum te facit 26:74 tunc coepit detestari et iurare quia non 26:74 Then began he to curse and to swear, saying, I know

not the man. And immediately the cock crew.

nouisset hominem et continuo gallus cantauit 26:75 et re- 26:75 And Peter remembered the word of Jesus, which

said unto him, Before the cock crow, thou shalt deny me thrice. And he went out, and wept bitterly.

cordatus est Petrus uerbi Iesu quod dixerat priusquam gallus cantet ter me negabis et egressus foras plorauit amare

27:1 mane autem facto consilium inierunt omnes princi- 27:1 When the morning was come, all the chief priests and

elders of the people took counsel against Jesus to put him to death:

pes sacerdotum et seniores populi aduersus Iesum ut eum morti traderent 27:2 et uinctum adduxerunt eum et tradi- 27:2 And when they had bound him, they led him away,

and delivered him to Pontius Pilate the governor.

derunt Pontio Pilato praesidi 27:3 tunc uidens Iudas qui 27:3 Then Judas, which had betrayed him, when he saw

that he was condemned, repented himself, and brought again the thirty pieces of silver to the chief priests and

eum tradidit quod damnatus esset paenitentia ductus rettu- elders, lit triginta argenteos principibus sacerdotum et senioribus 27:4 dicens peccaui tradens sanguinem iustum at illi dixerunt 27:4 Saying, I have sinned in that I have betrayed the in-

nocent blood. And they said, What is that to us? see thou to that.

quid ad nos tu uideris 27:5 et proiectis argenteis in templo

27:5 And he cast down the pieces of silver in the temple, and departed, and went and hanged himself.

recessit et abiens laqueo se suspendit 27:6 principes autem 27:6 And the chief priests took the silver pieces, and said,

It is not lawful for to put them into the treasury, because it is the price of blood.

sacerdotum acceptis argenteis dixerunt non licet mittere eos in corbanan quia pretium sanguinis est 27:7 consilio autem 27:7 And they took counsel, and bought with them the

potter’s field, to bury strangers in.

inito emerunt ex illis agrum figuli in sepulturam peregrinorum 27:8 propter hoc uocatus est ager ille Acheldemach 27:8 Wherefore that field was called, The field of blood,

unto this day.

ager sanguinis usque in hodiernum diem 27:9 tunc imple- 27:9 Then was fulfilled that which was spoken by Jeremy

the prophet, saying, And they took the thirty pieces of silver, the price of him that was valued, whom they of the

tum est quod dictum est per Hieremiam prophetam dicentem children of Israel did value; et acceperunt triginta argenteos pretium adpretiati quem adpretiauerunt a filiis Israhel 27:10 et dederunt eos in agrum 27:10 And gave them for the potter’s field, as the Lord

appointed me.

figuli sicut constituit mihi Dominus 27:11 Iesus autem stetit 27:11 And Jesus stood before the governor: and the gov-

ernor asked him, saying, Art thou the King of the Jews? And Jesus said unto him, Thou sayest.

ante praesidem et interrogauit eum praeses dicens tu es rex Iudaeorum dicit ei Iesus tu dicis 27:12 et cum accusaretur a 27:12 And when he was accused of the chief priests and

elders, he answered nothing.

principibus sacerdotum et senioribus nihil respondit 27:13 27:13 Then said Pilate unto him, Hearest thou not how

many things they witness against thee?

Biblia Sacra Vulgata

tunc dicit illi Pilatus non audis quanta aduersum te dicant

27:14 And he answered him to never a word; insomuch testimonia 27:14 et non respondit ei ad ullum uerbum ita

that the governor marvelled greatly.

27:15 Now at that feast the governor was wont to release ut miraretur praeses uehementer 27:15 per diem autem sol-

unto the people a prisoner, whom they would.

lemnem consueuerat praeses dimittere populo unum uinc-

27:16 And they had then a notable prisoner, called Barab- tum quem uoluissent 27:16 habebat autem tunc uinctum inbas. 27:17 Therefore when they were gathered together, Pilate signem qui dicebatur Barabbas 27:17 congregatis ergo ilsaid unto them, Whom will ye that I release unto you? Barabbas, or Jesus which is called Christ?

lis dixit Pilatus quem uultis dimittam uobis Barabban an Ie-

27:18 For he knew that for envy they had delivered him. sum qui dicitur Christus 27:18 sciebat enim quod per in-27:19 When he was set down on the judgment seat, his uidiam tradidissent eum 27:19 sedente autem illo pro tri-

wife sent unto him, saying, Have thou nothing to do with that just man: for I have suffered many things this day in

a dream because of him. bunali misit ad illum uxor eius dicens nihil tibi et iusto illi

27:20 But the chief priests and elders persuaded the mul- multa enim passa sum hodie per uisum propter eum 27:20 titude that they should ask Barabbas, and destroy Jesus.

princeps autem sacerdotum et seniores persuaserunt populis

27:21 The governor answered and said unto them, ut peterent Barabban Iesum uero perderent 27:21 respon-

Whether of the twain will ye that I release unto you? They said, Barabbas.

dens autem praeses ait illis quem uultis uobis de duobus di-

27:22 Pilate saith unto them, What shall I do then with mitti at illi dixerunt Barabban 27:22 dicit illis Pilatus quid

Jesus which is called Christ? They all say unto him, Let him be crucified.

igitur faciam de Iesu qui dicitur Christus 27:23 dicunt om-

27:23 And the governor said, Why, what evil hath he done? But they cried out the more, saying, Let him be crucified.

nes crucifigatur ait illis praeses quid enim mali fecit at illi

27:24 When Pilate saw that he could prevail nothing, but magis clamabant dicentes crucifigatur 27:24 uidens autem

that rather a tumult was made, he took water, and washed his hands before the multitude, saying, I am innocent of

the blood of this just person: see ye to it. Pilatus quia nihil proficeret sed magis tumultus fieret ac-

cepta aqua lauit manus coram populo dicens innocens ego

27:25 Then answered all the people, and said, His blood sum a sanguine iusti huius uos uideritis 27:25 et respon-

be on us, and on our children.

dens uniuersus populus dixit sanguis eius super nos et super

27:26 Then released he Barabbas unto them: and when he filios nostros 27:26 tunc dimisit illis Barabban Iesum autem

had scourged Jesus, he delivered him to be crucified.

27:27 Then the soldiers of the governor took Jesus into flagellatum tradidit eis ut crucifigeretur 27:27 tunc milites

the common hall, and gathered unto him the whole band of soldiers.

praesidis suscipientes Iesum in praetorio congregauerunt ad

27:28 And they stripped him, and put on him a scarlet eum uniuersam cohortem 27:28 et exuentes eum clamydem

robe.

27:29 And when they had platted a crown of thorns, they coccineam circumdederunt ei 27:29 et plectentes coronam

put it upon his head, and a reed in his right hand: and they bowed the knee before him, and mocked him, saying,

Hail, King of the Jews! de spinis posuerunt super caput eius et harundinem in de-

xtera eius et genu flexo ante eum inludebant dicentes haue

27:30 And they spit upon him, and took the reed, and rex Iudaeorum 27:30 et expuentes in eum acceperunt ha-

smote him on the head.

27:31 And after that they had mocked him, they took the rundinem et percutiebant caput eius 27:31 et postquam in-

robe off from him, and put his own raiment on him, and led him away to crucify him.

luserunt ei exuerunt eum clamydem et induerunt eum uesti-

27:32 And as they came out, they found a man of Cyrene, mentis eius et duxerunt eum ut crucifigerent 27:32 exeun-

Simon by name: him they compelled to bear his cross.

tes autem inuenerunt hominem cyreneum nomine Simonem hunc angariauerunt ut tolleret crucem eius 27:33 et uene- 27:33 And when they were come unto a place called Gol-

gotha, that is to say, a place of a skull,

runt in locum qui dicitur Golgotha quod est Caluariae lo-

cus 27:34 et dederunt ei uinum bibere cum felle mixtum 27:34 They gave him vinegar to drink mingled with gall:

and when he had tasted thereof, he would not drink.

et cum gustasset noluit bibere 27:35 postquam autem cru- 27:35 And they crucified him, and parted his garments,

casting lots: that it might be fulfilled which was spoken by the prophet, They parted my garments among them,

cifixerunt eum diuiserunt uestimenta eius sortem mittentes and upon my vesture did they cast lots. 27:36 et sedentes seruabant eum 27:37 et inposuerunt super 27:36 And sitting down they watched him there;

27:37 And set up over his head his accusation written, THIS IS JESUS THE KING OF THE JEWS.

caput eius causam ipsius scriptam hic est Iesus rex Iudaeo-

rum 27:38 tunc crucifixi sunt cum eo duo latrones unus a 27:38 Then were there two thieves crucified with him, one

on the right hand, and another on the left.

dextris et unus a sinistris 27:39 praetereuntes autem blas- 27:39 And they that passed by reviled him, wagging their

heads,

phemabant eum mouentes capita sua 27:40 et dicentes qui 27:40 And saying, Thou that destroyest the temple, and

buildest it in three days, save thyself. If thou be the Son of God, come down from the cross.

destruit templum et in triduo illud reaedificat salua temet ipsum si Filius Dei es descende de cruce 27:41 similiter 27:41 Likewise also the chief priests mocking him, with

the scribes and elders, said,

et principes sacerdotum inludentes cum scribis et senioribus dicentes 27:42 alios saluos fecit se ipsum non potest 27:42 He saved others; himself he cannot save. If he be

the King of Israel, let him now come down from the cross, and we will believe him.

saluum facere si rex Israhel est descendat nunc de cruce et credemus ei 27:43 confidet in Deo liberet nunc eum si uult 27:43 He trusted in God; let him deliver him now, if he

will have him: for he said, I am the Son of God.

dixit enim quia Dei Filius sum 27:44 id ipsum autem et la- 27:44 The thieves also, which were crucified with him,

cast the same in his teeth.

trones qui fixi erant cum eo inproperabant ei 27:45 a sexta 27:45 Now from the sixth hour there was darkness over all

the land unto the ninth hour.

autem hora tenebrae factae sunt super uniuersam terram usque ad horam nonam 27:46 et circa horam nonam clamauit 27:46 And about the ninth hour Jesus cried with a loud

voice, saying, Eli, Eli, lama sabachthani? that is to say, My God, my God, why hast thou forsaken me?

Iesus uoce magna dicens Heli Heli lema sabacthani hoc est

Deus meus Deus meus ut quid dereliquisti me 27:47 qui- 27:47 Some of them that stood there, when they heard that,

said, This man calleth for Elias.

dam autem illic stantes et audientes dicebant Heliam uocat iste 27:48 et continuo currens unus ex eis acceptam spon- 27:48 And straightway one of them ran, and took a spunge,

and filled it with vinegar, and put it on a reed, and gave him to drink.

giam impleuit aceto et inposuit harundini et dabat ei bibere 27:49 ceteri uero dicebant sine uideamus an ueniat Helias li- 27:49 The rest said, Let be, let us see whether Elias will

come to save him.

berans eum 27:50 Iesus autem iterum clamans uoce magna 27:50 Jesus, when he had cried again with a loud voice,

yielded up the ghost.

emisit spiritum 27:51 et ecce uelum templi scissum est in 27:51 And, behold, the veil of the temple was rent in twain

from the top to the bottom; and the earth did quake, and the rocks rent;

duas partes a summo usque deorsum et terra mota est et petrae scissae sunt 27:52 et monumenta aperta sunt et multa 27:52 And the graves were opened; and many bodies of

the saints which slept arose,

corpora sanctorum qui dormierant surrexerunt 27:53 et ex- 27:53 And came out of the graves after his resurrection,

and went into the holy city, and appeared unto many.

euntes de monumentis post resurrectionem eius uenerunt in

Biblia Sacra Vulgata

27:54 Now when the centurion, and they that were with sanctam ciuitatem et apparuerunt multis 27:54 centurio au-

him, watching Jesus, saw the earthquake, and those things that were done, they feared greatly, saying, Truly this was

the Son of God. tem et qui cum eo erant custodientes Iesum uiso terraemotu

et his quae fiebant timuerunt ualde dicentes uere Dei Filius

27:55 And many women were there beholding afar off, erat iste 27:55 erant autem ibi mulieres multae a longe quae

which followed Jesus from Galilee, ministering unto him:

27:56 Among which was Mary Magdalene, and Mary the secutae erant Iesum a Galilaea ministrantes ei 27:56 inter

mother of James and Joses, and the mother of Zebedees children.

quas erat Maria Magdalene et Maria Iacobi et Ioseph mater

27:57 When the even was come, there came a rich man of et mater filiorum Zebedaei 27:57 cum sero autem factum

Arimathaea, named Joseph, who also himself was Jesus’ disciple:

esset uenit quidam homo diues ab Arimathia nomine Ioseph

27:58 He went to Pilate, and begged the body of Jesus. qui et ipse discipulus erat Iesu 27:58 hic accessit ad Pilatum

Then Pilate commanded the body to be delivered.

27:59 And when Joseph had taken the body, he wrapped it et petiit corpus Iesu tunc Pilatus iussit reddi corpus 27:59 et

in a clean linen cloth,

27:60 And laid it in his own new tomb, which he had hewn accepto corpore Ioseph inuoluit illud sindone munda 27:60

out in the rock: and he rolled a great stone to the door of the sepulchre, and departed.

et posuit illud in monumento suo nouo quod exciderat in petra et aduoluit saxum magnum ad ostium monumenti et

27:61 And there was Mary Magdalene, and the other abiit 27:61 erat autem ibi Maria Magdalene et altera Maria

Mary, sitting over against the sepulchre.

27:62 Now the next day, that followed the day of the sedentes contra sepulchrum 27:62 altera autem die quae est

preparation, the chief priests and Pharisees came together unto Pilate,

post parasceuen conuenerunt principes sacerdotum et Pha-

27:63 Saying, Sir, we remember that that deceiver said, risaei ad Pilatum 27:63 dicentes domine recordati sumus

while he was yet alive, After three days I will rise again.

quia seductor ille dixit adhuc uiuens post tres dies resur-

27:64 Command therefore that the sepulchre be made sure gam 27:64 iube ergo custodiri sepulchrum usque in diem until the third day, lest his disciples come by night, and steal him away, and say unto the people, He is risen from

the dead: so the last error shall be worse than the first. tertium ne forte ueniant discipuli eius et furentur eum et di-

cant plebi surrexit a mortuis et erit nouissimus error peior

27:65 Pilate said unto them, Ye have a watch: go your priore 27:65 ait illis Pilatus habetis custodiam ite custodite

way, make it as sure as ye can.

27:66 So they went, and made the sepulchre sure, sealing sicut scitis 27:66 illi autem abeuntes munierunt sepulchrum

the stone, and setting a watch.

signantes lapidem cum custodibus

28:1 In the end of the sabbath, as it began to dawn toward 28:1 uespere autem sabbati quae lucescit in primam sab-

the first day of the week, came Mary Magdalene and the other Mary to see the sepulchre.

bati uenit Maria Magdalene et altera Maria uidere sepul-

28:2 And, behold, there was a great earthquake: for the chrum 28:2 et ecce terraemotus factus est magnus angelus

angel of the Lord descended from heaven, and came and rolled back the stone from the door, and sat upon it.

enim Domini descendit de caelo et accedens reuoluit lapi-

28:3 His countenance was like lightning, and his raiment dem et sedebat super eum 28:3 erat autem aspectus eius white as snow:

28:4 And for fear of him the keepers did shake, and be- sicut fulgur et uestimentum eius sicut nix 28:4 prae timore

came as dead men.

autem eius exterriti sunt custodes et facti sunt uelut mor-

28:5 And the angel answered and said unto the women, tui 28:5 respondens autem angelus dixit mulieribus nolite

Fear not ye: for I know that ye seek Jesus, which was crucified.

timere uos scio enim quod Iesum qui crucifixus est quaeri-

Euangelium Secundum Matthaeum

tis 28:6 non est hic surrexit enim sicut dixit uenite uidete 28:6 He is not here: for he is risen, as he said. Come, see

the place where the Lord lay.

locum ubi positus erat Dominus 28:7 et cito euntes dicite 28:7 And go quickly, and tell his disciples that he is risen from the dead; and, behold, he goeth before you into Galilee; there shall ye see him: lo, I have told you.

discipulis eius quia surrexit et ecce praecedit uos in Galilaeam ibi eum uidebitis ecce praedixi uobis 28:8 et exierunt 28:8 And they departed quickly from the sepulchre with

fear and great joy; and did run to bring his disciples word.

cito de monumento cum timore et magno gaudio currentes nuntiare discipulis eius 28:9 et ecce Iesus occurrit illis di- 28:9 And as they went to tell his disciples, behold, Jesus

met them, saying, All hail. And they came and held him by the feet, and worshipped him.

cens hauete illae autem accesserunt et tenuerunt pedes eius et adorauerunt eum 28:10 tunc ait illis Iesus nolite timere ite 28:10 Then said Jesus unto them, Be not afraid: go tell my

brethren that they go into Galilee, and there shall they see me.

nuntiate fratribus meis ut eant in Galilaeam ibi me uidebunt 28:11 quae cum abissent ecce quidam de custodibus uenerunt 28:11 Now when they were going, behold, some of the

watch came into the city, and shewed unto the chief priests all the things that were done.

in ciuitatem et nuntiauerunt principibus sacerdotum omnia quae facta fuerant 28:12 et congregati cum senioribus con- 28:12 And when they were assembled with the elders, and

had taken counsel, they gave large money unto the soldiers,

silio accepto pecuniam copiosam dederunt militibus 28:13

28:13 Saying, Say ye, His disciples came by night, and stole him away while we slept.

dicentes dicite quia discipuli eius nocte uenerunt et furati sunt eum nobis dormientibus 28:14 et si hoc auditum fuerit 28:14 And if this come to the governor’s ears, we will per-

suade him, and secure you.

a praeside nos suadebimus ei et securos uos faciemus 28:15 28:15 So they took the money, and did as they were taught:

and this saying is commonly reported among the Jews until this day.

at illi accepta pecunia fecerunt sicut erant docti et diuulgatum est uerbum istud apud Iudaeos usque in hodiernum diem 28:16 undecim autem discipuli abierunt in Galilaeam in mon- 28:16 Then the eleven disciples went away into Galilee,

into a mountain where Jesus had appointed them.

tem ubi constituerat illis Iesus 28:17 et uidentes eum adora- 28:17 And when they saw him, they worshipped him: but

some doubted.

uerunt quidam autem dubitauerunt 28:18 et accedens Iesus 28:18 And Jesus came and spake unto them, saying, All

power is given unto me in heaven and in earth.

locutus est eis dicens data est mihi omnis potestas in caelo et in terra 28:19 euntes ergo docete omnes gentes baptizantes 28:19 Go ye therefore, and teach all nations, baptizing

them in the name of the Father, and of the Son, and of the Holy Ghost:

eos in nomine Patris et Filii et Spiritus Sancti 28:20 docen-

28:20 Teaching them to observe all things whatsoever I have commanded you: and, lo, I am with you alway, even unto the end of the world. Amen.

tes eos seruare omnia quaecumque mandaui uobis et ecce ego uobiscum sum omnibus diebus usque ad consummationem saeculi

Biblia Sacra Vulgata

MCD

Euangelium Secundum Marcum

1:1 initium euangelii Iesu Christi Filii Dei 1:2 sicut scri- 1:1 The beginning of the gospel of Jesus Christ, the Son

of God;

1:2 As it is written in the prophets, Behold, I send my messenger before thy face, which shall prepare thy way before thee.

ptum est in Esaia propheta ecce mitto angelum meum ante faciem tuam qui praeparabit uiam tuam 1:3 uox clamantis 1:3 The voice of one crying in the wilderness, Prepare ye

the way of the Lord, make his paths straight.

in deserto parate uiam Domini rectas facite semitas eius 1:4 1:4 John did baptize in the wilderness, and preach the bap-

tism of repentance for the remission of sins.

fuit Iohannes in deserto baptizans et praedicans baptismum paenitentiae in remissionem peccatorum 1:5 et egredieba- 1:5 And there went out unto him all the land of Judaea,

and they of Jerusalem, and were all baptized of him in the river of Jordan, confessing their sins.

tur ad illum omnis Iudaeae regio et Hierosolymitae uniuersi et baptizabantur ab illo in Iordane flumine confitentes peccata sua 1:6 et erat Iohannes uestitus pilis cameli et zona 1:6 And John was clothed with camel’s hair, and with a

girdle of a skin about his loins; and he did eat locusts and wild honey;

pellicia circa lumbos eius et lucustas et mel siluestre edebat 1:7 et praedicabat dicens uenit fortior me post me cuius 1:7 And preached, saying, There cometh one mightier than

I after me, the latchet of whose shoes I am not worthy to stoop down and unloose.

non sum dignus procumbens soluere corrigiam calciamentorum eius 1:8 ego baptizaui uos aqua ille uero baptizabit uos 1:8 I indeed have baptized you with water: but he shall

baptize you with the Holy Ghost.

Spiritu Sancto 1:9 et factum est in diebus illis uenit Iesus a 1:9 And it came to pass in those days, that Jesus came from

Nazareth of Galilee, and was baptized of John in Jordan.

Nazareth Galilaeae et baptizatus est in Iordane ab Iohanne 1:10 et statim ascendens de aqua uidit apertos caelos et Spiri- 1:10 And straightway coming up out of the water, he saw

the heavens opened, and the Spirit like a dove descending upon him:

tum tamquam columbam descendentem et manentem in ipso 1:11 et uox facta est de caelis tu es Filius meus dilectus in te 1:11 And there came a voice from heaven, saying, Thou

art my beloved Son, in whom I am well pleased.

conplacui 1:12 et statim Spiritus expellit eum in desertum 1:12 And immediately the spirit driveth him into the

wilderness.

1:13 et erat in deserto quadraginta diebus et quadraginta noc- 1:13 And he was there in the wilderness forty days,

tempted of Satan; and was with the wild beasts; and the angels ministered unto him.

tibus et temptabatur a Satana eratque cum bestiis et angeli ministrabant illi 1:14 postquam autem traditus est Iohan- 1:14 Now after that John was put in prison, Jesus came

into Galilee, preaching the gospel of the kingdom of God,

nes uenit Iesus in Galilaeam praedicans euangelium regni Dei 1:15 et dicens quoniam impletum est tempus et adpro- 1:15 And saying, The time is fulfilled, and the kingdom of

God is at hand: repent ye, and believe the gospel.

pinquauit regnum Dei paenitemini et credite euangelio 1:16 1:16 Now as he walked by the sea of Galilee, he saw Si-

mon and Andrew his brother casting a net into the sea: for they were fishers.

et praeteriens secus mare Galilaeae uidit Simonem et Andream fratrem eius mittentes retia in mare erant enim piscatores 1:17 et dixit eis Iesus uenite post me et faciam uos fieri 1:17 And Jesus said unto them, Come ye after me, and I

will make you to become fishers of men.

piscatores hominum 1:18 et protinus relictis retibus secuti 1:18 And straightway they forsook their nets, and followed

him.

sunt eum 1:19 et progressus inde pusillum uidit Iacobum 1:19 And when he had gone a little farther thence, he saw

James the son of Zebedee, and John his brother, who also were in the ship mending their nets.

Zebedaei et Iohannem fratrem eius et ipsos in naui conponentes retia 1:20 et statim uocauit illos et relicto patre suo 1:20 And straightway he called them: and they left their

father Zebedee in the ship with the hired servants, and went after him.

Zebedaeo in naui cum mercennariis secuti sunt eum 1:21

1:21 And they went into Capernaum; and straightway on the sabbath day he entered into the synagogue, and taught.

Biblia Sacra Vulgata

et ingrediuntur Capharnaum et statim sabbatis ingressus sy-

1:22 And they were astonished at his doctrine: for he nagogam docebat eos 1:22 et stupebant super doctrina eius

taught them as one that had authority, and not as the scribes.

erat enim docens eos quasi potestatem habens et non sicut

1:23 And there was in their synagogue a man with an un- scribae 1:23 et erat in synagoga eorum homo in spiritu in-

clean spirit; and he cried out,

1:24 Saying, Let us alone; what have we to do with thee, mundo et exclamauit 1:24 dicens quid nobis et tibi Iesu Na-

thou Jesus of Nazareth? art thou come to destroy us? I know thee who thou art, the Holy One of God.

zarene uenisti perdere nos scio qui sis Sanctus Dei 1:25 et

1:25 And Jesus rebuked him, saying, Hold thy peace, and come out of him.

comminatus est ei Iesus dicens obmutesce et exi de homine

1:26 And when the unclean spirit had torn him, and cried 1:26 et discerpens eum spiritus inmundus et exclamans uoce

with a loud voice, he came out of him. 1:27 And they were all amazed, insomuch that they ques- magna exiuit ab eo 1:27 et mirati sunt omnes ita ut contioned among themselves, saying, What thing is this? what new doctrine is this? for with authority commandeth

he even the unclean spirits, and they do obey him. quirerent inter se dicentes quidnam est hoc quae doctrina

haec noua quia in potestate et spiritibus inmundis imperat et

1:28 And immediately his fame spread abroad throughout oboediunt ei 1:28 et processit rumor eius statim in omnem

all the region round about Galilee.

1:29 And forthwith, when they were come out of the syn- regionem Galilaeae 1:29 et protinus egredientes de syna-

agogue, they entered into the house of Simon and Andrew, with James and John.

goga uenerunt in domum Simonis et Andreae cum Iacobo

1:30 But Simon’s wife’s mother lay sick of a fever, and et Iohanne 1:30 decumbebat autem socrus Simonis febri-

anon they tell him of her.

1:31 And he came and took her by the hand, and lifted her citans et statim dicunt ei de illa 1:31 et accedens eleuauit

up; and immediately the fever left her, and she ministered unto them.

eam adprehensa manu eius et continuo dimisit eam febris et

1:32 And at even, when the sun did set, they brought unto ministrabat eis 1:32 uespere autem facto cum occidisset sol

him all that were diseased, and them that were possessed with devils.

adferebant ad eum omnes male habentes et daemonia haben-

1:33 And all the city was gathered together at the door. tes 1:33 et erat omnis ciuitas congregata ad ianuam 1:34 et

1:34 And he healed many that were sick of divers diseases, and cast out many devils; and suffered not the devils to speak, because they knew him.

curauit multos qui uexabantur uariis languoribus et daemonia multa eiciebat et non sinebat loqui ea quoniam sciebant

1:35 And in the morning, rising up a great while before eum 1:35 et diluculo ualde surgens egressus abiit in deser-

day, he went out, and departed into a solitary place, and there prayed.

tum locum ibique orabat 1:36 et persecutus est eum Simon

1:36 And Simon and they that were with him followed after him.

1:37 And when they had found him, they said unto him, et qui cum illo erant 1:37 et cum inuenissent eum dixerunt

All men seek for thee.

1:38 And he said unto them, Let us go into the next towns, ei quia omnes quaerunt te 1:38 et ait illis eamus in proxi-

that I may preach there also: for therefore came I forth.

mos uicos et ciuitates ut et ibi praedicem ad hoc enim ueni

1:39 And he preached in their synagogues throughout all 1:39 et erat praedicans in synagogis eorum et omni Galilaea

Galilee, and cast out devils.

1:40 And there came a leper to him, beseeching him, and et daemonia eiciens 1:40 et uenit ad eum leprosus depre-

kneeling down to him, and saying unto him, If thou wilt, thou canst make me clean.

cans eum et genu flexo dixit si uis potes me mundare 1:41

1:41 And Jesus, moved with compassion, put forth his hand, and touched him, and saith unto him, I will; be thou clean.

Iesus autem misertus eius extendit manum suam et tangens

1:42 And as soon as he had spoken, immediately the lep- eum ait illi uolo mundare 1:42 et cum dixisset statim dis-

rosy departed from him, and he was cleansed. 1:43 And he straitly charged him, and forthwith sent him cessit ab eo lepra et mundatus est 1:43 et comminatus ei away;

Euangelium Secundum Marcum

statim eiecit illum 1:44 et dicit ei uide nemini dixeris sed 1:44 And saith unto him, See thou say nothing to any man:

but go thy way, shew thyself to the priest, and offer for thy cleansing those things which Moses commanded, for

uade ostende te principi sacerdotum et offer pro emunda- a testimony unto them. tione tua quae praecepit Moses in testimonium illis 1:45 at 1:45 But he went out, and began to publish it much, and

to blaze abroad the matter, insomuch that Jesus could no more openly enter into the city, but was without in desert

ille egressus coepit praedicare et diffamare sermonem ita ut places: and they came to him from every quarter. iam non posset manifeste in ciuitatem introire sed foris in desertis locis esse et conueniebant ad eum undique

2:1 et iterum intrauit Capharnaum post dies 2:2 et audi- 2:1 And again he entered into Capernaum after some days;

and it was noised that he was in the house.

2:2 And straightway many were gathered together, insomuch that there was no room to receive them, no, not so much as about the door: and he preached the word unto them.

tum est quod in domo esset et conuenerunt multi ita ut non caperet neque ad ianuam et loquebatur eis uerbum 2:3 et

2:3 And they come unto him, bringing one sick of the palsy, which was borne of four.

uenerunt ferentes ad eum paralyticum qui a quattuor portabatur 2:4 et cum non possent offerre eum illi prae turba 2:4 And when they could not come nigh unto him for the

press, they uncovered the roof where he was: and when they had broken it up, they let down the bed wherein the

nudauerunt tectum ubi erat et patefacientes submiserunt gra- sick of the palsy lay. battum in quo paralyticus iacebat 2:5 cum uidisset autem 2:5 When Jesus saw their faith, he said unto the sick of the

palsy, Son, thy sins be forgiven thee.

Iesus fidem illorum ait paralytico fili dimittuntur tibi peccata 2:6 erant autem illic quidam de scribis sedentes et co- 2:6 But there was certain of the scribes sitting there, and

reasoning in their hearts,

gitantes in cordibus suis 2:7 quid hic sic loquitur blasphe- 2:7 Why doth this man thus speak blasphemies? who can

forgive sins but God only?

mat quis potest dimittere peccata nisi solus Deus 2:8 quo 2:8 And immediately when Jesus perceived in his spirit

that they so reasoned within themselves, he said unto them, Why reason ye these things in your hearts?

statim cognito Iesus spiritu suo quia sic cogitarent intra se dicit illis quid ista cogitatis in cordibus uestris 2:9 quid est 2:9 Whether is it easier to say to the sick of the palsy, Thy

sins be forgiven thee; or to say, Arise, and take up thy bed, and walk?

facilius dicere paralytico dimittuntur tibi peccata an dicere surge et tolle grabattum tuum et ambula 2:10 ut autem scia- 2:10 But that ye may know that the Son of man hath power

on earth to forgive sins, (he saith to the sick of the palsy,)

tis quia potestatem habet Filius hominis in terra dimittendi peccata ait paralytico 2:11 tibi dico surge tolle grabattum 2:11 I say unto thee, Arise, and take up thy bed, and go

thy way into thine house.

tuum et uade in domum tuam 2:12 et statim ille surrexit 2:12 And immediately he arose, took up the bed, and went forth before them all; insomuch that they were all amazed, and glorified God, saying, We never saw it on this fashion.

et sublato grabatto abiit coram omnibus ita ut admirarentur omnes et honorificarent Deum dicentes quia numquam sic uidimus 2:13 et egressus est rursus ad mare omnisque turba 2:13 And he went forth again by the sea side; and all the

multitude resorted unto him, and he taught them.

ueniebat ad eum et docebat eos 2:14 et cum praeteriret ui- 2:14 And as he passed by, he saw Levi the son of Alphaeus

sitting at the receipt of custom, and said unto him, Follow me. And he arose and followed him.

dit Leuin Alphei sedentem ad teloneum et ait illi sequere

me et surgens secutus est eum 2:15 et factum est cum ac- 2:15 And it came to pass, that, as Jesus sat at meat in his house, many publicans and sinners sat also together with Jesus and his disciples: for there were many, and they fol-

cumberet in domo illius multi publicani et peccatores simul lowed him. discumbebant cum Iesu et discipulis eius erant enim multi qui et sequebantur eum 2:16 et scribae et Pharisaei uiden- 2:16 And when the scribes and Pharisees saw him eat with

publicans and sinners, they said unto his disciples, How is it that he eateth and drinketh with publicans and sinners?

Biblia Sacra Vulgata

tes quia manducaret cum peccatoribus et publicanis dicebant discipulis eius quare cum publicanis et peccatoribus mandu-

2:17 When Jesus heard it, he saith unto them, They that are cat et bibit magister uester 2:17 hoc audito Iesus ait illis

whole have no need of the physician, but they that are sick: I came not to call the righteous, but sinners to repentance.

non necesse habent sani medicum sed qui male habent non

2:18 And the disciples of John and of the Pharisees used enim ueni uocare iustos sed peccatores 2:18 et erant dis-

to fast: and they come and say unto him, Why do the disciples of John and of the Pharisees fast, but thy disciples

fast not? cipuli Iohannis et Pharisaei ieiunantes et ueniunt et dicunt

illi cur discipuli Iohannis et Pharisaeorum ieiunant tui autem

2:19 And Jesus said unto them, Can the children of the discipuli non ieiunant 2:19 et ait illis Iesus numquid pos-

bridechamber fast, while the bridegroom is with them? as long as they have the bridegroom with them, they cannot

fast. sunt filii nuptiarum quamdiu sponsus cum illis est ieiunare

quanto tempore habent secum sponsum non possunt ieiu-

2:20 But the days will come, when the bridegroom shall nare 2:20 uenient autem dies cum auferetur ab eis sponsus

be taken away from them, and then shall they fast in those days.

et tunc ieiunabunt in illa die 2:21 nemo adsumentum panni

2:21 No man also seweth a piece of new cloth on an old garment: else the new piece that filled it up taketh away from the old, and the rent is made worse.

rudis adsuit uestimento ueteri alioquin aufert supplementum

2:22 And no man putteth new wine into old bottles: else nouum a ueteri et maior scissura fit 2:22 et nemo mittit ui-

the new wine doth burst the bottles, and the wine is spilled, and the bottles will be marred: but new wine must be put

into new bottles. num nouellum in utres ueteres alioquin disrumpet uinum ut-

res et uinum effunditur et utres peribunt sed uinum nouum

2:23 And it came to pass, that he went through the corn in utres nouos mitti debet 2:23 et factum est iterum cum

fields on the sabbath day; and his disciples began, as they went, to pluck the ears of corn.

sabbatis ambularet per sata et discipuli eius coeperunt prae-

2:24 And the Pharisees said unto him, Behold, why do gredi et uellere spicas 2:24 Pharisaei autem dicebant ei ecce

they on the sabbath day that which is not lawful?

2:25 And he said unto them, Have ye never read what quid faciunt sabbatis quod non licet 2:25 et ait illis num-

David did, when he had need, and was an hungred, he, and they that were with him?

quam legistis quid fecerit Dauid quando necessitatem habuit

2:26 How he went into the house of God in the days of et esuriit ipse et qui cum eo erant 2:26 quomodo introiit

Abiathar the high priest, and did eat the shewbread, which is not lawful to eat but for the priests, and gave also to

them which were with him? in domum Dei sub Abiathar principe sacerdotum et panes

propositionis manducauit quos non licet manducare nisi sa-

2:27 And he said unto them, The sabbath was made for cerdotibus et dedit eis qui cum eo erant 2:27 et dicebat eis

man, and not man for the sabbath:

sabbatum propter hominem factum est et non homo prop-

2:28 Therefore the Son of man is Lord also of the sabbath. ter sabbatum 2:28 itaque dominus est Filius hominis etiam sabbati

3:1 And he entered again into the synagogue; and there 3:1 et introiuit iterum synagogam et erat ibi homo habens

was a man there which had a withered hand.

3:2 And they watched him, whether he would heal him on manum aridam 3:2 et obseruabant eum si sabbatis curaret

the sabbath day; that they might accuse him.

3:3 And he saith unto the man which had the withered ut accusarent illum 3:3 et ait homini habenti manum ari-

hand, Stand forth.

3:4 And he saith unto them, Is it lawful to do good on the dam surge in medium 3:4 et dicit eis licet sabbatis bene

sabbath days, or to do evil? to save life, or to kill? But they held their peace.

facere an male animam saluam facere an perdere at illi ta-

Euangelium Secundum Marcum

cebant 3:5 et circumspiciens eos cum ira contristatus su- 3:5 And when he had looked round about on them with

anger, being grieved for the hardness of their hearts, he saith unto the man, Stretch forth thine hand. And he stretched it out: and his hand was restored whole as the other.

per caecitatem cordis eorum dicit homini extende manum tuam et extendit et restituta est manus illi 3:6 exeuntes au- 3:6 And the Pharisees went forth, and straightway took

counsel with the Herodians against him, how they might destroy him.

tem statim Pharisaei cum Herodianis consilium faciebant ad-

uersus eum quomodo eum perderent 3:7 et Iesus cum di- 3:7 But Jesus withdrew himself with his disciples to the sea: and a great multitude from Galilee followed him, and from Judaea,

scipulis suis secessit ad mare et multa turba a Galilaea et Iudaea secuta est eum 3:8 et ab Hierosolymis et ab Idu- 3:8 And from Jerusalem, and from Idumaea, and from be-

yond Jordan; and they about Tyre and Sidon, a great multitude, when they had heard what great things he did, came

mea et trans Iordanen et qui circa Tyrum et Sidonem mul- unto him. titudo magna audientes quae faciebat uenerunt ad eum 3:9 3:9 And he spake to his disciples, that a small ship should

wait on him because of the multitude, lest they should throng him.

et dixit discipulis suis ut nauicula sibi deseruiret propter turbam ne conprimerent eum 3:10 multos enim sanabat ita ut 3:10 For he had healed many; insomuch that they pressed

upon him for to touch him, as many as had plagues.

inruerent in eum ut illum tangerent quotquot habebant plagas 3:11 et spiritus inmundi cum illum uidebant procide- 3:11 And unclean spirits, when they saw him, fell down

before him, and cried, saying, Thou art the Son of God.

bant ei et clamabant dicentes 3:12 tu es Filius Dei et ue- 3:12 And he straitly charged them that they should not

make him known.

hementer comminabatur eis ne manifestarent illum 3:13 et 3:13 And he goeth up into a mountain, and calleth unto

him whom he would: and they came unto him.

ascendens in montem uocauit ad se quos uoluit ipse et uenerunt ad eum 3:14 et fecit ut essent duodecim cum illo et 3:14 And he ordained twelve, that they should be with

him, and that he might send them forth to preach,

ut mitteret eos praedicare 3:15 et dedit illis potestatem cu- 3:15 And to have power to heal sicknesses, and to cast out

devils:

randi infirmitates et eiciendi daemonia 3:16 et inposuit Si- 3:16 And Simon he surnamed Peter; moni nomen Petrus 3:17 et Iacobum Zebedaei et Iohannem 3:17 And James the son of Zebedee, and John the brother

of James; and he surnamed them Boanerges, which is, The sons of thunder:

fratrem Iacobi et inposuit eis nomina Boanerges quod est Filii tonitrui 3:18 et Andream et Philippum et Bartholomeum 3:18 And Andrew, and Philip, and Bartholomew, and

Matthew, and Thomas, and James the son of Alphaeus, and Thaddaeus, and Simon the Canaanite,

et Mattheum et Thomam et Iacobum Alphei et Thaddeum et Simonem Cananeum 3:19 et Iudam Scarioth qui et tradidit 3:19 And Judas Iscariot, which also betrayed him: and

they went into an house.

illum 3:20 et ueniunt ad domum et conuenit iterum turba ita 3:20 And the multitude cometh together again, so that they

could not so much as eat bread.

ut non possent neque panem manducare 3:21 et cum audis- 3:21 And when his friends heard of it, they went out to lay

hold on him: for they said, He is beside himself.

sent sui exierunt tenere eum dicebant enim quoniam in furorem uersus est 3:22 et scribae qui ab Hierosolymis descen- 3:22 And the scribes which came down from Jerusalem

said, He hath Beelzebub, and by the prince of the devils casteth he out devils.

derant dicebant quoniam Beelzebub habet et quia in principe daemonum eicit daemonia 3:23 et conuocatis eis in pa- 3:23 And he called them unto him, and said unto them in

parables, How can Satan cast out Satan?

rabolis dicebat illis quomodo potest Satanas Satanan eicere

3:24 et si regnum in se diuidatur non potest stare regnum 3:24 And if a kingdom be divided against itself, that king-

dom cannot stand.

illud 3:25 et si domus super semet ipsam dispertiatur non 3:25 And if a house be divided against itself, that house

cannot stand.

Biblia Sacra Vulgata

3:26 And if Satan rise up against himself, and be divided, poterit domus illa stare 3:26 et si Satanas consurrexit in se-

he cannot stand, but hath an end.

met ipsum dispertitus est et non potest stare sed finem habet

3:27 No man can enter into a strong man’s house, and spoil 3:27 nemo potest uasa fortis ingressus in domum diripere nisi

his goods, except he will first bind the strong man; and then he will spoil his house.

prius fortem alliget et tunc domum eius diripiet 3:28 amen

3:28 Verily I say unto you, All sins shall be forgiven unto the sons of men, and blasphemies wherewith soever they shall blaspheme:

dico uobis quoniam omnia dimittentur filiis hominum pec-

3:29 But he that shall blaspheme against the Holy Ghost cata et blasphemiae quibus blasphemauerint 3:29 qui autem

hath never forgiveness, but is in danger of eternal damnation.

blasphemauerit in Spiritum Sanctum non habet remissionem

3:30 Because they said, He hath an unclean spirit. in aeternum sed reus erit aeterni delicti 3:30 quoniam dice-3:31 There came then his brethren and his mother, and, bant spiritum inmundum habet 3:31 et ueniunt mater eius et

standing without, sent unto him, calling him.

3:32 And the multitude sat about him, and they said unto fratres et foris stantes miserunt ad eum uocantes eum 3:32

him, Behold, thy mother and thy brethren without seek for thee.

et sedebat circa eum turba et dicunt ei ecce mater tua et fra-

3:33 And he answered them, saying, Who is my mother, tres tui foris quaerunt te 3:33 et respondens eis ait quae est

or my brethren?

3:34 And he looked round about on them which sat about mater mea et fratres mei 3:34 et circumspiciens eos qui in

him, and said, Behold my mother and my brethren!

3:35 For whosoever shall do the will of God, the same is circuitu eius sedebant ait ecce mater mea et fratres mei 3:35

my brother, and my sister, and mother.

qui enim fecerit uoluntatem Dei hic frater meus et soror mea et mater est

4:1 And he began again to teach by the sea side: and there 4:1 et iterum coepit docere ad mare et congregata est ad was gathered unto him a great multitude, so that he entered into a ship, and sat in the sea; and the whole multitude was

by the sea on the land. eum turba multa ita ut in nauem ascendens sederet in mari 4:2 And he taught them many things by parables, and said et omnis turba circa mare super terram erat 4:2 et docebat

unto them in his doctrine,

4:3 Hearken; Behold, there went out a sower to sow: eos in parabolis multa et dicebat illis in doctrina sua 4:3 4:4 And it came to pass, as he sowed, some fell by the way audite ecce exiit seminans ad seminandum 4:4 et dum se-

side, and the fowls of the air came and devoured it up.

minat aliud cecidit circa uiam et uenerunt uolucres et co-

4:5 And some fell on stony ground, where it had not much mederunt illud 4:5 aliud uero cecidit super petrosa ubi non

earth; and immediately it sprang up, because it had no depth of earth:

habuit terram multam et statim exortum est quoniam non ha-

4:6 But when the sun was up, it was scorched; and because bebat altitudinem terrae 4:6 et quando exortus est sol exae-

it had no root, it withered away.

4:7 And some fell among thorns, and the thorns grew up, stuauit et eo quod non haberet radicem exaruit 4:7 et aliud

and choked it, and it yielded no fruit.

cecidit in spinas et ascenderunt spinae et offocauerunt illud

4:8 And other fell on good ground, and did yield fruit that et fructum non dedit 4:8 et aliud cecidit in terram bonam et

sprang up and increased; and brought forth, some thirty, and some sixty, and some an hundred.

dabat fructum ascendentem et crescentem et adferebat unum

4:9 And he said unto them, He that hath ears to hear, let triginta et unum sexaginta et unum centum 4:9 et dicebat

him hear.

4:10 And when he was alone, they that were about him qui habet aures audiendi audiat 4:10 et cum esset singu-

with the twelve asked of him the parable.

laris interrogauerunt eum hii qui cum eo erant cum duode-

4:11 And he said unto them, Unto you it is given to know cim parabolas 4:11 et dicebat eis uobis datum est myste-

the mystery of the kingdom of God: but unto them that are without, all these things are done in parables:

Euangelium Secundum Marcum

rium regni Dei illis autem qui foris sunt in parabolis omnia fiunt 4:12 ut uidentes uideant et non uideant et audientes 4:12 That seeing they may see, and not perceive; and hear-

ing they may hear, and not understand; lest at any time they should be converted, and their sins should be forgiven

audiant et non intellegant nequando conuertantur et dimit- them. tantur eis peccata 4:13 et ait illis nescitis parabolam hanc 4:13 And he said unto them, Know ye not this parable?

and how then will ye know all parables?

et quomodo omnes parabolas cognoscetis 4:14 qui seminat 4:14 The sower soweth the word. uerbum seminat 4:15 hii autem sunt qui circa uiam ubi se- 4:15 And these are they by the way side, where the word

is sown; but when they have heard, Satan cometh immediately, and taketh away the word that was sown in their

minatur uerbum et cum audierint confestim uenit Satanas et hearts. aufert uerbum quod seminatum est in corda eorum 4:16 et 4:16 And these are they likewise which are sown on stony

ground; who, when they have heard the word, immediately receive it with gladness;

hii sunt similiter qui super petrosa seminantur qui cum audierint uerbum statim cum gaudio accipiunt illud 4:17 et 4:17 And have no root in themselves, and so endure but for

a time: afterward, when affliction or persecution ariseth for the word’s sake, immediately they are offended.

non habent radicem in se sed temporales sunt deinde orta tribulatione et persecutione propter uerbum confestim scandalizantur 4:18 et alii sunt qui in spinis seminantur hii sunt 4:18 And these are they which are sown among thorns;

such as hear the word,

qui uerbum audiunt 4:19 et aerumnae saeculi et deceptio 4:19 And the cares of this world, and the deceitfulness of riches, and the lusts of other things entering in, choke the word, and it becometh unfruitful.

diuitiarum et circa reliqua concupiscentiae introeuntes suffocant uerbum et sine fructu efficitur 4:20 et hii sunt qui 4:20 And these are they which are sown on good ground;

such as hear the word, and receive it, and bring forth fruit, some thirtyfold, some sixty, and some an hundred.

super terram bonam seminati sunt qui audiunt uerbum et suscipiunt et fructificant unum triginta et unum sexaginta et unum centum 4:21 et dicebat illis numquid uenit lucerna ut 4:21 And he said unto them, Is a candle brought to be put

under a bushel, or under a bed? and not to be set on a candlestick?

sub modio ponatur aut sub lecto nonne ut super candelabrum ponatur 4:22 non enim est aliquid absconditum quod 4:22 For there is nothing hid, which shall not be mani-

fested; neither was any thing kept secret, but that it should come abroad.

non manifestetur nec factum est occultum sed ut in palam ueniat 4:23 si quis habet aures audiendi audiat 4:24 et di- 4:23 If any man have ears to hear, let him hear.

4:24 And he said unto them, Take heed what ye hear: with what measure ye mete, it shall be measured to you: and unto you that hear shall more be given.

cebat illis uidete quid audiatis in qua mensura mensi fueritis remetietur uobis et adicietur uobis 4:25 qui enim habet 4:25 For he that hath, to him shall be given: and he that

hath not, from him shall be taken even that which he hath.

dabitur illi et qui non habet etiam quod habet auferetur ab illo 4:26 et dicebat sic est regnum Dei quemadmodum si 4:26 And he said, So is the kingdom of God, as if a man

should cast seed into the ground;

homo iaciat sementem in terram 4:27 et dormiat et exsur- 4:27 And should sleep, and rise night and day, and the seed

should spring and grow up, he knoweth not how.

gat nocte ac die et semen germinet et increscat dum nescit ille 4:28 ultro enim terra fructificat primum herbam deinde 4:28 For the earth bringeth forth fruit of herself; first the

blade, then the ear, after that the full corn in the ear.

spicam deinde plenum frumentum in spica 4:29 et cum se 4:29 But when the fruit is brought forth, immediately he

putteth in the sickle, because the harvest is come.

produxerit fructus statim mittit falcem quoniam adest messis 4:30 et dicebat cui adsimilabimus regnum Dei aut cui para- 4:30 And he said, Whereunto shall we liken the kingdom

of God? or with what comparison shall we compare it?

Biblia Sacra Vulgata

4:31 It is like a grain of mustard seed, which, when it is bolae conparabimus illud 4:31 sicut granum sinapis quod

sown in the earth, is less than all the seeds that be in the earth:

cum seminatum fuerit in terra minus est omnibus seminibus

4:32 But when it is sown, it groweth up, and becometh quae sunt in terra 4:32 et cum seminatum fuerit ascendit et

greater than all herbs, and shooteth out great branches; so that the fowls of the air may lodge under the shadow of it.

fit maius omnibus holeribus et facit ramos magnos ita ut pos-

4:33 And with many such parables spake he the word unto sint sub umbra eius aues caeli habitare 4:33 et talibus multis

them, as they were able to hear it.

4:34 But without a parable spake he not unto them: and parabolis loquebatur eis uerbum prout poterant audire 4:34

when they were alone, he expounded all things to his disciples.

sine parabola autem non loquebatur eis seorsum autem disci-

4:35 And the same day, when the even was come, he saith pulis suis disserebat omnia 4:35 et ait illis illa die cum sero

unto them, Let us pass over unto the other side.

4:36 And when they had sent away the multitude, they esset factum transeamus contra 4:36 et dimittentes turbam

took him even as he was in the ship. And there were also with him other little ships.

adsumunt eum ita ut erat in naui et aliae naues erant cum

4:37 And there arose a great storm of wind, and the waves illo 4:37 et facta est procella magna uenti et fluctus mit-

beat into the ship, so that it was now full.

4:38 And he was in the hinder part of the ship, asleep on tebat in nauem ita ut impleretur nauis 4:38 et erat ipse in

a pillow: and they awake him, and say unto him, Master, carest thou not that we perish?

puppi supra ceruical dormiens et excitant eum et dicunt ei

4:39 And he arose, and rebuked the wind, and said unto magister non ad te pertinet quia perimus 4:39 et exsurgens

the sea, Peace, be still. And the wind ceased, and there was a great calm.

comminatus est uento et dixit mari tace obmutesce et ces-

4:40 And he said unto them, Why are ye so fearful? how sauit uentus et facta est tranquillitas magna 4:40 et ait illis

is it that ye have no faith? And they feared exceedingly, and said one to another, What manner of man is this, that

even the wind and the sea obey him? quid timidi estis necdum habetis fidem et timuerunt magno timore et dicebant ad alterutrum quis putas est iste quia et uentus et mare oboediunt ei

5:1 And they came over unto the other side of the sea, into 5:1 et uenerunt trans fretum maris in regionem Geraseno-

the country of the Gadarenes.

5:2 And when he was come out of the ship, immediately rum 5:2 et exeunti ei de naui statim occurrit ei de monu-

there met him out of the tombs a man with an unclean spirit,

mentis homo in spiritu inmundo 5:3 qui domicilium habe-

5:3 Who had his dwelling among the tombs; and no man could bind him, no, not with chains:

bat in monumentis et neque catenis iam quisquam eum pote-

5:4 Because that he had been often bound with fetters and rat ligare 5:4 quoniam saepe conpedibus et catenis uinctus

chains, and the chains had been plucked asunder by him, and the fetters broken in pieces: neither could any man

tame him. disrupisset catenas et conpedes comminuisset et nemo pote-

5:5 And always, night and day, he was in the mountains, rat eum domare 5:5 et semper nocte ac die in monumentis and in the tombs, crying, and cutting himself with stones.

5:6 But when he saw Jesus afar off, he ran and worshipped et in montibus erat clamans et concidens se lapidibus 5:6

him,

5:7 And cried with a loud voice, and said, What have I uidens autem Iesum a longe cucurrit et adorauit eum 5:7

to do with thee, Jesus, thou Son of the most high God? I adjure thee by God, that thou torment me not.

et clamans uoce magna dicit quid mihi et tibi Iesu Fili Dei

5:8 For he said unto him, Come out of the man, thou un- summi adiuro te per Deum ne me torqueas 5:8 dicebat enim

clean spirit.

5:9 And he asked him, What is thy name? And he an- illi exi spiritus inmunde ab homine 5:9 et interrogabat eum

swered, saying, My name is Legion: for we are many.

quod tibi nomen est et dicit ei Legio nomen mihi est quia

5:10 And he besought him much that he would not send multi sumus 5:10 et deprecabatur eum multum ne se expel-

them away out of the country.

Euangelium Secundum Marcum

leret extra regionem 5:11 erat autem ibi circa montem grex 5:11 Now there was there nigh unto the mountains a great

herd of swine feeding.

porcorum magnus pascens 5:12 et deprecabantur eum spi- 5:12 And all the devils besought him, saying, Send us into

the swine, that we may enter into them.

ritus dicentes mitte nos in porcos ut in eos introeamus 5:13 5:13 And forthwith Jesus gave them leave. And the un-

clean spirits went out, and entered into the swine: and the herd ran violently down a steep place into the sea, (they

et concessit eis statim Iesus et exeuntes spiritus inmundi in- were about two thousand;) and were choked in the sea. troierunt in porcos et magno impetu grex praecipitatus est in mare ad duo milia et suffocati sunt in mare 5:14 qui autem 5:14 And they that fed the swine fled, and told it in the

city, and in the country. And they went out to see what it was that was done.

pascebant eos fugerunt et nuntiauerunt in ciuitatem et in agros et egressi sunt uidere quid esset facti 5:15 et ueniunt ad 5:15 And they come to Jesus, and see him that was pos-

sessed with the devil, and had the legion, sitting, and clothed, and in his right mind: and they were afraid.

Iesum et uident illum qui a daemonio uexabatur sedentem uestitum et sanae mentis et timuerunt 5:16 et narrauerunt 5:16 And they that saw it told them how it befell to him

that was possessed with the devil, and also concerning the swine.

illis qui uiderant qualiter factum esset ei qui daemonium habuerat et de porcis 5:17 et rogare eum coeperunt ut discede- 5:17 And they began to pray him to depart out of their

coasts.

ret de finibus eorum 5:18 cumque ascenderet nauem coepit 5:18 And when he was come into the ship, he that had

been possessed with the devil prayed him that he might be with him.

illum deprecari qui daemonio uexatus fuerat ut esset cum illo 5:19 et non admisit eum sed ait illi uade in domum tuam ad 5:19 Howbeit Jesus suffered him not, but saith unto him,

Go home to thy friends, and tell them how great things the Lord hath done for thee, and hath had compassion on thee.

tuos et adnuntia illis quanta tibi Dominus fecerit et misertus sit tui 5:20 et abiit et coepit praedicare in Decapoli quanta 5:20 And he departed, and began to publish in Decapolis

how great things Jesus had done for him: and all men did marvel.

sibi fecisset Iesus et omnes mirabantur 5:21 et cum trans-

5:21 And when Jesus was passed over again by ship unto the other side, much people gathered unto him: and he was nigh unto the sea.

cendisset Iesus in naui rursus trans fretum conuenit turba multa ad illum et erat circa mare 5:22 et uenit quidam de ar- 5:22 And, behold, there cometh one of the rulers of the

synagogue, Jairus by name; and when he saw him, he fell at his feet,

chisynagogis nomine Iairus et uidens eum procidit ad pedes

eius 5:23 et deprecabatur eum multum dicens quoniam filia 5:23 And besought him greatly, saying, My little daughter lieth at the point of death: I pray thee, come and lay thy hands on her, that she may be healed; and she shall live.

mea in extremis est ueni inpone manus super eam ut salua sit et uiuat 5:24 et abiit cum illo et sequebatur eum turba multa 5:24 And Jesus went with him; and much people followed

him, and thronged him.

et conprimebant illum 5:25 et mulier quae erat in profluuio 5:25 And a certain woman, which had an issue of blood

twelve years,

sanguinis annis duodecim 5:26 et fuerat multa perpessa a 5:26 And had suffered many things of many physicians,

and had spent all that she had, and was nothing bettered, but rather grew worse,

conpluribus medicis et erogauerat omnia sua nec quicquam profecerat sed magis deterius habebat 5:27 cum audisset de 5:27 When she had heard of Jesus, came in the press be-

hind, and touched his garment.

Iesu uenit in turba retro et tetigit uestimentum eius 5:28 di- 5:28 For she said, If I may touch but his clothes, I shall be

whole.

cebat enim quia si uel uestimentum eius tetigero salua ero 5:29 et confestim siccatus est fons sanguinis eius et sensit 5:29 And straightway the fountain of her blood was dried

up; and she felt in her body that she was healed of that plague.

corpore quod sanata esset a plaga 5:30 et statim Iesus cog-

5:30 And Jesus, immediately knowing in himself that virtue had gone out of him, turned him about in the press, and said, Who touched my clothes?

noscens in semet ipso uirtutem quae exierat de eo conuersus

Biblia Sacra Vulgata

5:31 And his disciples said unto him, Thou seest the mul- ad turbam aiebat quis tetigit uestimenta mea 5:31 et dice-

titude thronging thee, and sayest thou, Who touched me?

bant ei discipuli sui uides turbam conprimentem te et dicis

5:32 And he looked round about to see her that had done quis me tetigit 5:32 et circumspiciebat uidere eam quae hoc

this thing.

5:33 But the woman fearing and trembling, knowing what fecerat 5:33 mulier autem timens et tremens sciens quod

was done in her, came and fell down before him, and told him all the truth.

factum esset in se uenit et procidit ante eum et dixit ei om-

5:34 And he said unto her, Daughter, thy faith hath made nem ueritatem 5:34 ille autem dixit ei filia fides tua te sal-

thee whole; go in peace, and be whole of thy plague.

5:35 While he yet spake, there came from the ruler of uam fecit uade in pace et esto sana a plaga tua 5:35 ad-

the synagogue’s house certain which said, Thy daughter is dead: why troublest thou the Master any further?

huc eo loquente ueniunt ab archisynagogo dicentes quia filia

5:36 As soon as Jesus heard the word that was spoken, he tua mortua est quid ultra uexas magistrum 5:36 Iesus autem

saith unto the ruler of the synagogue, Be not afraid, only believe.

uerbo quod dicebatur audito ait archisynagogo noli timere

5:37 And he suffered no man to follow him, save Peter, tantummodo crede 5:37 et non admisit quemquam sequi se

and James, and John the brother of James.

5:38 And he cometh to the house of the ruler of the syn- nisi Petrum et Iacobum et Iohannem fratrem Iacobi 5:38 et

agogue, and seeth the tumult, and them that wept and wailed greatly.

ueniunt in domum archisynagogi et uidet tumultum et flen-

5:39 And when he was come in, he saith unto them, Why tes et heiulantes multum 5:39 et ingressus ait eis quid tur-

make ye this ado, and weep? the damsel is not dead, but sleepeth.

bamini et ploratis puella non est mortua sed dormit 5:40

5:40 And they laughed him to scorn. But when he had put them all out, he taketh the father and the mother of the damsel, and them that were with him, and entereth in where the damsel was lying.

et inridebant eum ipse uero eiectis omnibus adsumit patrem et matrem puellae et qui secum erant et ingreditur ubi erat

5:41 And he took the damsel by the hand, and said unto puella iacens 5:41 et tenens manum puellae ait illi talitha

her, Talitha cumi; which is, being interpreted, Damsel, I say unto thee, arise.

cumi quod est interpretatum puella tibi dico surge 5:42 et

5:42 And straightway the damsel arose, and walked; for she was of the age of twelve years. And they were astonished with a great astonishment.

confestim surrexit puella et ambulabat erat autem annorum

5:43 And he charged them straitly that no man should duodecim et obstipuerunt stupore maximo 5:43 et praecepit

know it; and commanded that something should be given her to eat.

illis uehementer ut nemo id sciret et dixit dari illi manducare

6:1 And he went out from thence, and came into his own 6:1 et egressus inde abiit in patriam suam et sequebantur

country; and his disciples follow him.

6:2 And when the sabbath day was come, he began to illum discipuli sui 6:2 et facto sabbato coepit in synagoga

teach in the synagogue: and many hearing him were astonished, saying, From whence hath this man these things? and what wisdom is this which is given unto him, that even such mighty works are wrought by his hands?

docere et multi audientes admirabantur in doctrina eius dicentes unde huic haec omnia et quae est sapientia quae data

6:3 Is not this the carpenter, the son of Mary, the brother est illi et uirtutes tales quae per manus eius efficiuntur 6:3

of James, and Joses, and of Juda, and Simon? and are not his sisters here with us? And they were offended at him.

nonne iste est faber filius Mariae frater Iacobi et Ioseph et Iudae et Simonis nonne et sorores eius hic nobiscum sunt

6:4 But Jesus, said unto them, A prophet is not without et scandalizabantur in illo 6:4 et dicebat eis Iesus quia non

honour, but in his own country, and among his own kin, and in his own house.

est propheta sine honore nisi in patria sua et in cognatione

6:5 And he could there do no mighty work, save that he sua et in domo sua 6:5 et non poterat ibi uirtutem ullam fa-

laid his hands upon a few sick folk, and healed them.

6:6 And he marvelled because of their unbelief. And he cere nisi paucos infirmos inpositis manibus curauit 6:6 et

went round about the villages, teaching.

Euangelium Secundum Marcum

mirabatur propter incredulitatem eorum 6:7 et circumibat 6:7 And he called unto him the twelve, and began to send

them forth by two and two; and gave them power over unclean spirits;

castella in circuitu docens et conuocauit duodecim et coepit eos mittere binos et dabat illis potestatem spirituum inmundorum 6:8 et praecepit eis ne quid tollerent in uia nisi uir- 6:8 And commanded them that they should take nothing

for their journey, save a staff only; no scrip, no bread, no money in their purse:

gam tantum non peram non panem neque in zona aes 6:9

6:9 But be shod with sandals; and not put on two coats.

sed calciatos sandaliis et ne induerentur duabus tunicis 6:10 6:10 And he said unto them, In what place soever ye enter

into an house, there abide till ye depart from that place.

et dicebat eis quocumque introieritis in domum illic manete donec exeatis inde 6:11 et quicumque non receperint uos 6:11 And whosoever shall not receive you, nor hear you,

when ye depart thence, shake off the dust under your feet for a testimony against them. Verily I say unto you, It shall be more tolerable for Sodom and Gomorrha in the day of judgment, than for that city.

nec audierint uos exeuntes inde excutite puluerem de pedibus uestris in testimonium illis 6:12 et exeuntes praedica- 6:12 And they went out, and preached that men should

repent.

bant ut paenitentiam agerent 6:13 et daemonia multa eicie- 6:13 And they cast out many devils, and anointed with oil

many that were sick, and healed them.

bant et unguebant oleo multos aegrotos et sanabant 6:14 et 6:14 And king Herod heard of him; (for his name was

spread abroad:) and he said, That John the Baptist was risen from the dead, and therefore mighty works do shew

audiuit Herodes rex manifestum enim factum est nomen eius forth themselves in him. et dicebat quia Iohannes Baptista resurrexit a mortuis et propterea inoperantur uirtutes in illo 6:15 alii autem dicebant 6:15 Others said, That it is Elias. And others said, That it

is a prophet, or as one of the prophets.

quia Helias est alii uero dicebant propheta est quasi unus ex prophetis 6:16 quo audito Herodes ait quem ego decollaui 6:16 But when Herod heard thereof, he said, It is John,

whom I beheaded: he is risen from the dead.

Iohannem hic a mortuis resurrexit 6:17 ipse enim Herodes 6:17 For Herod himself had sent forth and laid hold upon

John, and bound him in prison for Herodias’ sake, his brother Philip’s wife: for he had married her.

misit ac tenuit Iohannem et uinxit eum in carcere propter Herodiadem uxorem Philippi fratris sui quia duxerat eam 6:18 6:18 For John had said unto Herod, It is not lawful for thee

to have thy brother’s wife.

dicebat enim Iohannes Herodi non licet tibi habere uxorem fratris tui 6:19 Herodias autem insidiabatur illi et uolebat 6:19 Therefore Herodias had a quarrel against him, and

would have killed him; but she could not:

occidere eum nec poterat 6:20 Herodes enim metuebat Io- 6:20 For Herod feared John, knowing that he was a just

man and an holy, and observed him; and when he heard him, he did many things, and heard him gladly.

hannem sciens eum uirum iustum et sanctum et custodiebat eum et audito eo multa faciebat et libenter eum audiebat 6:21 et cum dies oportunus accidisset Herodes natalis sui ce- 6:21 And when a convenient day was come, that Herod on

his birthday made a supper to his lords, high captains, and chief estates of Galilee;

nam fecit principibus et tribunis et primis Galilaeae 6:22

6:22 And when the daughter of the said Herodias came in, and danced, and pleased Herod and them that sat with him, the king said unto the damsel, Ask of me whatsoever thou wilt, and I will give it thee.

cumque introisset filia ipsius Herodiadis et saltasset et placuisset Herodi simulque recumbentibus rex ait puellae pete a me quod uis et dabo tibi 6:23 et iurauit illi quia quicquid pe- 6:23 And he sware unto her, Whatsoever thou shalt ask of

me, I will give it thee, unto the half of my kingdom.

tieris dabo tibi licet dimidium regni mei 6:24 quae cum ex- 6:24 And she went forth, and said unto her mother, What

shall I ask? And she said, The head of John the Baptist.

isset dixit matri suae quid petam et illa dixit caput Iohannis Baptistae 6:25 cumque introisset statim cum festinatione ad 6:25 And she came in straightway with haste unto the

king, and asked, saying, I will that thou give me by and by in a charger the head of John the Baptist.

Biblia Sacra Vulgata

regem petiuit dicens uolo ut protinus des mihi in disco ca-

6:26 And the king was exceeding sorry; yet for his oath’s put Iohannis Baptistae 6:26 et contristatus rex propter iu-

sake, and for their sakes which sat with him, he would not reject her.

siurandum et propter simul recumbentes noluit eam contris-

6:27 And immediately the king sent an executioner, and tare 6:27 sed misso speculatore praecepit adferri caput eius

commanded his head to be brought: and he went and beheaded him in the prison,

in disco et decollauit eum in carcere 6:28 et adtulit caput

6:28 And brought his head in a charger, and gave it to the damsel: and the damsel gave it to her mother.

eius in disco et dedit illud puellae et puella dedit matri suae

6:29 And when his disciples heard of it, they came and 6:29 quo audito discipuli eius uenerunt et tulerunt corpus eius

took up his corpse, and laid it in a tomb.

6:30 And the apostles gathered themselves together unto et posuerunt illud in monumento 6:30 et conuenientes apo-

Jesus, and told him all things, both what they had done, and what they had taught.

stoli ad Iesum renuntiauerunt illi omnia quae egerant et do-

6:31 And he said unto them, Come ye yourselves apart cuerant 6:31 et ait illis uenite seorsum in desertum locum

into a desert place, and rest a while: for there were many coming and going, and they had no leisure so much as to

eat. et requiescite pusillum erant enim qui ueniebant et rediibant 6:32 And they departed into a desert place by ship pri- multi et nec manducandi spatium habebant 6:32 et ascen-

vately.

6:33 And the people saw them departing, and many knew dentes in naui abierunt in desertum locum seorsum 6:33 et

him, and ran afoot thither out of all cities, and outwent them, and came together unto him.

uiderunt eos abeuntes et cognouerunt multi et pedestre et de omnibus ciuitatibus concurrerunt illuc et praeuenerunt eos

6:34 And Jesus, when he came out, saw much people, and 6:34 et exiens uidit multam turbam Iesus et misertus est super

was moved with compassion toward them, because they were as sheep not having a shepherd: and he began to

teach them many things. eos quia erant sicut oues non habentes pastorem et coepit do-6:35 And when the day was now far spent, his disciples cere illos multa 6:35 et cum iam hora multa fieret accesse-

came unto him, and said, This is a desert place, and now the time is far passed:

runt discipuli eius dicentes desertus est locus hic et iam hora

6:36 Send them away, that they may go into the country praeteriuit 6:36 dimitte illos ut euntes in proximas uillas et

round about, and into the villages, and buy themselves bread: for they have nothing to eat.

uicos emant sibi cibos quos manducent 6:37 et respondens

6:37 He answered and said unto them, Give ye them to eat. And they say unto him, Shall we go and buy two hundred pennyworth of bread, and give them to eat?

ait illis date illis manducare et dixerunt ei euntes emamus

6:38 He saith unto them, How many loaves have ye? go denariis ducentis panes et dabimus eis manducare 6:38 et

and see. And when they knew, they say, Five, and two fishes.

dicit eis quot panes habetis ite et uidete et cum cognouis-

6:39 And he commanded them to make all sit down by sent dicunt quinque et duos pisces 6:39 et praecepit illis ut

companies upon the green grass.

accumbere facerent omnes secundum contubernia super ui-

6:40 And they sat down in ranks, by hundreds, and by ride faenum 6:40 et discubuerunt in partes per centenos et

fifties.

6:41 And when he had taken the five loaves and the two per quinquagenos 6:41 et acceptis quinque panibus et duo-

fishes, he looked up to heaven, and blessed, and brake the loaves, and gave them to his disciples to set before them;

and the two fishes divided he among them all. bus piscibus intuens in caelum benedixit et fregit panes et

dedit discipulis suis ut ponerent ante eos et duos pisces diui-

6:42 And they did all eat, and were filled. sit omnibus 6:42 et manducauerunt omnes et saturati sunt 6:43 And they took up twelve baskets full of the frag- 6:43 et sustulerunt reliquias fragmentorum duodecim cofinos

ments, and of the fishes.

6:44 And they that did eat of the loaves were about five plenos et de piscibus 6:44 erant autem qui manducauerunt

thousand men.

Euangelium Secundum Marcum

quinque milia uirorum 6:45 et statim coegit discipulos suos 6:45 And straightway he constrained his disciples to get

into the ship, and to go to the other side before unto Bethsaida, while he sent away the people.

ascendere nauem ut praecederent eum trans fretum ad Bethsaidam dum ipse dimitteret populum 6:46 et cum dimisis- 6:46 And when he had sent them away, he departed into a

mountain to pray.

set eos abiit in montem orare 6:47 et cum sero esset erat 6:47 And when even was come, the ship was in the midst

of the sea, and he alone on the land.

nauis in medio mari et ipse solus in terra 6:48 et uidens eos 6:48 And he saw them toiling in rowing; for the wind

was contrary unto them: and about the fourth watch of the night he cometh unto them, walking upon the sea, and

laborantes in remigando erat enim uentus contrarius eis et would have passed by them. circa quartam uigiliam noctis uenit ad eos ambulans super mare et uolebat praeterire eos 6:49 at illi ut uiderunt eum 6:49 But when they saw him walking upon the sea, they

supposed it had been a spirit, and cried out:

ambulantem super mare putauerunt fantasma esse et exclamauerunt 6:50 omnes enim eum uiderunt et conturbati sunt 6:50 For they all saw him, and were troubled. And im-

mediately he talked with them, and saith unto them, Be of good cheer: it is I; be not afraid.

et statim locutus est cum eis et dixit illis confidite ego sum nolite timere 6:51 et ascendit ad illos in nauem et cessauit 6:51 And he went up unto them into the ship; and the wind

ceased: and they were sore amazed in themselves beyond measure, and wondered.

uentus et plus magis intra se stupebant 6:52 non enim intel-

6:52 For they considered not the miracle of the loaves: for their heart was hardened.

lexerant de panibus erat enim cor illorum obcaecatum 6:53 6:53 And when they had passed over, they came into the

land of Gennesaret, and drew to the shore.

et cum transfretassent peruenerunt in terram Gennesareth et adplicuerunt 6:54 cumque egressi essent de naui continuo 6:54 And when they were come out of the ship, straight-

way they knew him,

cognouerunt eum 6:55 et percurrentes uniuersam regionem 6:55 And ran through that whole region round about, and

began to carry about in beds those that were sick, where they heard he was.

illam coeperunt in grabattis eos qui se male habebant cir-

cumferre ubi audiebant eum esse 6:56 et quocumque intro- 6:56 And whithersoever he entered, into villages, or cities, or country, they laid the sick in the streets, and besought him that they might touch if it were but the border of his

ibat in uicos uel in uillas aut ciuitates in plateis ponebant garment: and as many as touched him were made whole. infirmos et deprecabantur eum ut uel fimbriam uestimenti eius tangerent et quotquot tangebant eum salui fiebant

7:1 et conueniunt ad eum Pharisaei et quidam de scribis 7:1 Then came together unto him the Pharisees, and cer-

tain of the scribes, which came from Jerusalem.

uenientes ab Hierosolymis 7:2 et cum uidissent quosdam 7:2 And when they saw some of his disciples eat bread

with defiled, that is to say, with unwashen, hands, they found fault.

ex discipulis eius communibus manibus id est non lotis manducare panes uituperauerunt 7:3 Pharisaei enim et omnes 7:3 For the Pharisees, and all the Jews, except they wash

their hands oft, eat not, holding the tradition of the elders.

Iudaei nisi crebro lauent manus non manducant tenentes traditionem seniorum 7:4 et a foro nisi baptizentur non co- 7:4 And when they come from the market, except they

wash, they eat not. And many other things there be, which they have received to hold, as the washing of cups, and

medunt et alia multa sunt quae tradita sunt illis seruare ba- pots, brasen vessels, and of tables. ptismata calicum et urceorum et aeramentorum et lectorum 7:5 et interrogant eum Pharisaei et scribae quare discipuli tui 7:5 Then the Pharisees and scribes asked him, Why walk

not thy disciples according to the tradition of the elders, but eat bread with unwashen hands?

non ambulant iuxta traditionem seniorum sed communibus manibus manducant panem 7:6 at ille respondens dixit eis 7:6 He answered and said unto them, Well hath Esaias

prophesied of you hypocrites, as it is written, This people honoureth me with their lips, but their heart is far from me.

Biblia Sacra Vulgata

bene prophetauit Esaias de uobis hypocritis sicut scriptum est populus hic labiis me honorat cor autem eorum longe

7:7 Howbeit in vain do they worship me, teaching for doc- est a me 7:7 in uanum autem me colunt docentes doctrinas trines the commandments of men.

7:8 For laying aside the commandment of God, ye hold praecepta hominum 7:8 relinquentes enim mandatum Dei

the tradition of men, as the washing of pots and cups: and many other such like things ye do.

tenetis traditionem hominum baptismata urceorum et cali-

7:9 And he said unto them, Full well ye reject the com- cum et alia similia his facitis multa 7:9 et dicebat illis bene

mandment of God, that ye may keep your own tradition.

irritum facitis praeceptum Dei ut traditionem uestram ser-

7:10 For Moses said, Honour thy father and thy mother; uetis 7:10 Moses enim dixit honora patrem tuum et matand, Whoso curseth father or mother, let him die the death:

rem tuam et qui maledixerit patri aut matri morte moriatur

7:11 But ye say, If a man shall say to his father or mother, 7:11 uos autem dicitis si dixerit homo patri aut matri corban

It is Corban, that is to say, a gift, by whatsoever thou mightest be profited by me; he shall be free.

quod est donum quodcumque ex me tibi profuerit 7:12 et

7:12 And ye suffer him no more to do ought for his father or his mother;

ultra non dimittitis eum quicquam facere patri suo aut matri

7:13 Making the word of God of none effect through your 7:13 rescindentes uerbum Dei per traditionem uestram quam

tradition, which ye have delivered: and many such like things do ye.

tradidistis et similia huiusmodi multa facitis 7:14 et aduo-

7:14 And when he had called all the people unto him, he said unto them, Hearken unto me every one of you, and understand:

cans iterum turbam dicebat illis audite me omnes et intel-

7:15 There is nothing from without a man, that entering legite 7:15 nihil est extra hominem introiens in eum quod

into him can defile him: but the things which come out of him, those are they that defile the man.

possit eum coinquinare sed quae de homine procedunt illa

7:16 If any man have ears to hear, let him hear. sunt quae communicant hominem 7:16 si quis habet aures 7:17 And when he was entered into the house from the audiendi audiat 7:17 et cum introisset in domum a turba

people, his disciples asked him concerning the parable.

7:18 And he saith unto them, Are ye so without under- interrogabant eum discipuli eius parabolam 7:18 et ait illis

standing also? Do ye not perceive, that whatsoever thing from without entereth into the man, it cannot defile him;

sic et uos inprudentes estis non intellegitis quia omne extrinsecus introiens in hominem non potest eum communicare

7:19 Because it entereth not into his heart, but into the 7:19 quia non introit in cor eius sed in uentrem et in secessum

belly, and goeth out into the draught, purging all meats?

7:20 And he said, That which cometh out of the man, that exit purgans omnes escas 7:20 dicebat autem quoniam quae

defileth the man. 7:21 For from within, out of the heart of men, proceed evil de homine exeunt illa communicant hominem 7:21 ab intus thoughts, adulteries, fornications, murders,

enim de corde hominum cogitationes malae procedunt adul-

7:22 Thefts, covetousness, wickedness, deceit, lascivious- teria fornicationes homicidia 7:22 furta auaritiae nequitiae

ness, an evil eye, blasphemy, pride, foolishness:

dolus inpudicitia oculus malus blasphemia superbia stultitia

7:23 All these evil things come from within, and defile the 7:23 omnia haec mala ab intus procedunt et communicant ho-

man.

7:24 And from thence he arose, and went into the borders minem 7:24 et inde surgens abiit in fines Tyri et Sidonis et

of Tyre and Sidon, and entered into an house, and would have no man know it: but he could not be hid.

ingressus domum neminem uoluit scire et non potuit latere

7:25 For a certain woman, whose young daughter had an 7:25 mulier enim statim ut audiuit de eo cuius habebat filia

unclean spirit, heard of him, and came and fell at his feet:

7:26 The woman was a Greek, a Syrophenician by nation; spiritum inmundum intrauit et procidit ad pedes eius 7:26

and she besought him that he would cast forth the devil out of her daughter.

Euangelium Secundum Marcum

erat autem mulier gentilis Syrophoenissa genere et rogabat

eum ut daemonium eiceret de filia eius 7:27 qui dixit illi 7:27 But Jesus said unto her, Let the children first be filled: for it is not meet to take the children’s bread, and to cast it unto the dogs.

sine prius saturari filios non est enim bonum sumere panem filiorum et mittere canibus 7:28 at illa respondit et dicit ei 7:28 And she answered and said unto him, Yes, Lord: yet

the dogs under the table eat of the children’s crumbs.

utique Domine nam et catelli sub mensa comedunt de micis puerorum 7:29 et ait illi propter hunc sermonem uade 7:29 And he said unto her, For this saying go thy way; the

devil is gone out of thy daughter.

exiit daemonium de filia tua 7:30 et cum abisset domum 7:30 And when she was come to her house, she found the

devil gone out, and her daughter laid upon the bed.

suam inuenit puellam iacentem supra lectum et daemonium exisse 7:31 et iterum exiens de finibus Tyri uenit per Sido- 7:31 And again, departing from the coasts of Tyre and

Sidon, he came unto the sea of Galilee, through the midst of the coasts of Decapolis.

nem ad mare Galilaeae inter medios fines Decapoleos 7:32

7:32 And they bring unto him one that was deaf, and had an impediment in his speech; and they beseech him to put his hand upon him.

et adducunt ei surdum et mutum et deprecantur eum ut inponat illi manum 7:33 et adprehendens eum de turba seorsum 7:33 And he took him aside from the multitude, and put his

fingers into his ears, and he spit, and touched his tongue;

misit digitos suos in auriculas et expuens tetigit linguam eius 7:34 et suspiciens in caelum ingemuit et ait illi eppheta quod 7:34 And looking up to heaven, he sighed, and saith unto

him, Ephphatha, that is, Be opened.

est adaperire 7:35 et statim apertae sunt aures eius et solu- 7:35 And straightway his ears were opened, and the string

of his tongue was loosed, and he spake plain.

tum est uinculum linguae eius et loquebatur recte 7:36 et 7:36 And he charged them that they should tell no man:

but the more he charged them, so much the more a great deal they published it;

praecepit illis ne cui dicerent quanto autem eis praecipiebat tanto magis plus praedicabant 7:37 et eo amplius admi- 7:37 And were beyond measure astonished, saying, He

hath done all things well: he maketh both the deaf to hear, and the dumb to speak.

rabantur dicentes bene omnia fecit et surdos facit audire et mutos loqui

8:1 in illis diebus iterum cum turba multa esset nec ha- 8:1 In those days the multitude being very great, and having nothing to eat, Jesus called his disciples unto him, and saith unto them,

berent quod manducarent conuocatis discipulis ait illis 8:2

8:2 I have compassion on the multitude, because they have now been with me three days, and have nothing to eat:

misereor super turba quia ecce iam triduo sustinent me nec habent quod manducent 8:3 et si dimisero eos ieiunos in 8:3 And if I send them away fasting to their own houses,

they will faint by the way: for divers of them came from far.

domum suam deficient in uia quidam enim ex eis de longe uenerunt 8:4 et responderunt ei discipuli sui unde istos po- 8:4 And his disciples answered him, From whence can a

man satisfy these men with bread here in the wilderness?

terit quis hic saturare panibus in solitudine 8:5 et interroga- 8:5 And he asked them, How many loaves have ye? And

they said, Seven.

uit eos quot panes habetis qui dixerunt septem 8:6 et prae- 8:6 And he commanded the people to sit down on the

ground: and he took the seven loaves, and gave thanks, and brake, and gave to his disciples to set before them;

cepit turbae discumbere supra terram et accipiens septem pa- and they did set them before the people. nes gratias agens fregit et dabat discipulis suis ut adponerent et adposuerunt turbae 8:7 et habebant pisciculos paucos et 8:7 And they had a few small fishes: and he blessed, and

commanded to set them also before them.

ipsos benedixit et iussit adponi 8:8 et manducauerunt et sa- 8:8 So they did eat, and were filled: and they took up of

the broken meat that was left seven baskets.

turati sunt et sustulerunt quod superauerat de fragmentis sep-8:9 And they that had eaten were about four thousand: and tem sportas 8:9 erant autem qui manducauerunt quasi quat-

he sent them away. 8:10 And straightway he entered into a ship with his dis- tuor milia et dimisit eos 8:10 et statim ascendens nauem ciples, and came into the parts of Dalmanutha.

8:11 And the Pharisees came forth, and began to question cum discipulis suis uenit in partes Dalmanutha 8:11 et exie-

with him, seeking of him a sign from heaven, tempting him.

runt Pharisaei et coeperunt conquirere cum eo quaerentes ab

8:12 And he sighed deeply in his spirit, and saith, Why illo signum de caelo temptantes eum 8:12 et ingemescens

doth this generation seek after a sign? verily I say unto you, There shall no sign be given unto this generation.

spiritu ait quid generatio ista quaerit signum amen dico uobis

8:13 And he left them, and entering into the ship again si dabitur generationi isti signum 8:13 et dimittens eos as-

departed to the other side.

8:14 Now the disciples had forgotten to take bread, neither cendens iterum abiit trans fretum 8:14 et obliti sunt sumere

had they in the ship with them more than one loaf.

8:15 And he charged them, saying, Take heed, beware of panes et nisi unum panem non habebant secum in naui 8:15

the leaven of the Pharisees, and of the leaven of Herod.

et praecipiebat eis dicens uidete cauete a fermento Pharisae-

8:16 And they reasoned among themselves, saying, It is orum et fermento Herodis 8:16 et cogitabant ad alterutrum

because we have no bread.

8:17 And when Jesus knew it, he saith unto them, Why dicentes quia panes non habemus 8:17 quo cognito Iesus

reason ye, because ye have no bread? perceive ye not yet, neither understand? have ye your heart yet hardened?

ait illis quid cogitatis quia panes non habetis nondum cognoscitis nec intellegitis adhuc caecatum habetis cor uestrum

8:18 Having eyes, see ye not? and having ears, hear ye 8:18 oculos habentes non uidetis et aures habentes non au-

not? and do ye not remember?

8:19 When I brake the five loaves among five thousand, ditis nec recordamini 8:19 quando quinque panes fregi in

how many baskets full of fragments took ye up? They say unto him, Twelve.

quinque milia et quot cofinos fragmentorum plenos sustu-

8:20 And when the seven among four thousand, how many listis dicunt ei duodecim 8:20 quando et septem panes in

baskets full of fragments took ye up? And they said, Seven.

quattuor milia quot sportas fragmentorum tulistis et dicunt

8:21 And he said unto them, How is it that ye do not un- ei septem 8:21 et dicebat eis quomodo nondum intellegitis

derstand?

8:22 And he cometh to Bethsaida; and they bring a blind 8:22 et ueniunt Bethsaida et adducunt ei caecum et rogabant

man unto him, and besought him to touch him.

8:23 And he took the blind man by the hand, and led him eum ut illum tangeret 8:23 et adprehendens manum caeci

out of the town; and when he had spit on his eyes, and put his hands upon him, he asked him if he saw ought.

eduxit eum extra uicum et expuens in oculos eius inposi-

8:24 And he looked up, and said, I see men as trees, walk- tis manibus suis interrogauit eum si aliquid uideret 8:24 et

ing.

8:25 After that he put his hands again upon his eyes, and aspiciens ait uideo homines uelut arbores ambulantes 8:25

made him look up: and he was restored, and saw every man clearly.

deinde iterum inposuit manus super oculos eius et coepit ui-

8:26 And he sent him away to his house, saying, Neither dere et restitutus est ita ut uideret clare omnia 8:26 et misit

go into the town, nor tell it to any in the town.

illum in domum suam dicens uade in domum tuam et si in

8:27 And Jesus went out, and his disciples, into the towns uicum introieris nemini dixeris 8:27 et egressus est Iesus

of Caesarea Philippi: and by the way he asked his disciples, saying unto them, Whom do men say that I am?

et discipuli eius in castella Caesareae Philippi et in uia interrogabat discipulos suos dicens eis quem me dicunt esse

8:28 And they answered, John the Baptist; but some say, homines 8:28 qui responderunt illi dicentes Iohannem Ba-

Elias; and others, One of the prophets.

8:29 And he saith unto them, But whom say ye that I ptistam alii Heliam alii uero quasi unum de prophetis 8:29

am? And Peter answereth and saith unto him, Thou art the Christ.

Euangelium Secundum Marcum

tunc dicit illis uos uero quem me dicitis esse respondens Petrus ait ei tu es Christus 8:30 et comminatus est eis ne cui 8:30 And he charged them that they should tell no man of

him.

dicerent de illo 8:31 et coepit docere illos quoniam opor- 8:31 And he began to teach them, that the Son of man

must suffer many things, and be rejected of the elders, and of the chief priests, and scribes, and be killed, and after

tet Filium hominis multa pati et reprobari a senioribus et a three days rise again. summis sacerdotibus et scribis et occidi et post tres dies resurgere 8:32 et palam uerbum loquebatur et adprehendens 8:32 And he spake that saying openly. And Peter took

him, and began to rebuke him.

eum Petrus coepit increpare eum 8:33 qui conuersus et ui- 8:33 But when he had turned about and looked on his dis-

ciples, he rebuked Peter, saying, Get thee behind me, Satan: for thou savourest not the things that be of God, but

dens discipulos suos comminatus est Petro dicens uade re- the things that be of men. tro me Satana quoniam non sapis quae Dei sunt sed quae sunt hominum 8:34 et conuocata turba cum discipulis suis 8:34 And when he had called the people unto him with his

disciples also, he said unto them, Whosoever will come after me, let him deny himself, and take up his cross, and

dixit eis si quis uult post me sequi deneget se ipsum et tollat follow me.

crucem suam et sequatur me 8:35 qui enim uoluerit ani- 8:35 For whosoever will save his life shall lose it; but whosoever shall lose his life for my sake and the gospel’s, the same shall save it.

mam suam saluam facere perdet eam qui autem perdiderit animam suam propter me et euangelium saluam eam faciet 8:36 quid enim prodeeit homini si lucretur mundum totum et 8:36 For what shall it profit a man, if he shall gain the

whole world, and lose his own soul?

detrimentum faciat animae suae 8:37 aut quid dabit homo 8:37 Or what shall a man give in exchange for his soul? commutationem pro anima sua 8:38 qui enim me confusus 8:38 Whosoever therefore shall be ashamed of me and of

my words in this adulterous and sinful generation; of him also shall the Son of man be ashamed, when he cometh in

fuerit et mea uerba in generatione ista adultera et peccatrice the glory of his Father with the holy angels. et Filius hominis confundetur eum cum uenerit in gloria Patris sui cum angelis sanctis 8:39 et dicebat illis amen dico 8:39 And he said unto them, Verily I say unto you, That

there be some of them that stand here, which shall not taste of death, till they have seen the kingdom of God come with

uobis quia sunt quidam de hic stantibus qui non gustabunt power. mortem donec uideant regnum Dei ueniens in uirtute

9:1 et post dies sex adsumit Iesus Petrum et Iacobum et 9:1 And after six days Jesus taketh with him Peter, and

James, and John, and leadeth them up into an high mountain apart by themselves: and he was transfigured before

Iohannem et ducit illos in montem excelsum seorsum so- them. los et transfiguratus est coram ipsis 9:2 et uestimenta eius 9:2 And his raiment became shining, exceeding white as

snow; so as no fuller on earth can white them.

facta sunt splendentia candida nimis uelut nix qualia fullo super terram non potest candida facere 9:3 et apparuit illis 9:3 And there appeared unto them Elias with Moses: and

they were talking with Jesus.

Helias cum Mose et erant loquentes cum Iesu 9:4 et res- 9:4 And Peter answered and said to Jesus, Master, it is

good for us to be here: and let us make three tabernacles; one for thee, and one for Moses, and one for Elias.

pondens Petrus ait Iesu rabbi bonum est hic nos esse et faciamus tria tabernacula tibi unum et Mosi unum et Heliae unum 9:5 non enim sciebat quid diceret erant enim timore 9:5 For he wist not what to say; for they were sore afraid. exterriti 9:6 et facta est nubes obumbrans eos et uenit uox 9:6 And there was a cloud that overshadowed them: and

a voice came out of the cloud, saying, This is my beloved Son: hear him.

de nube dicens hic est Filius meus carissimus audite illum

Biblia Sacra Vulgata

9:7 And suddenly, when they had looked round about, they 9:7 et statim circumspicientes neminem amplius uiderunt nisi

saw no man any more, save Jesus only with themselves.

9:8 And as they came down from the mountain, he charged Iesum tantum secum 9:8 et descendentibus illis de monte

them that they should tell no man what things they had seen, till the Son of man were risen from the dead.

praecepit illis ne cui quae uidissent narrarent nisi cum Filius

9:9 And they kept that saying with themselves, question- hominis a mortuis resurrexerit 9:9 et uerbum continuerunt

ing one with another what the rising from the dead should mean.

apud se conquirentes quid esset cum a mortuis resurrexerit

9:10 And they asked him, saying, Why say the scribes that 9:10 et interrogabant eum dicentes quid ergo dicunt Pharisaei

Elias must first come?

9:11 And he answered and told them, Elias verily cometh et scribae quia Heliam oporteat uenire primum 9:11 qui re-

first, and restoreth all things; and how it is written of the Son of man, that he must suffer many things, and be set at

nought. spondens ait illis Helias cum uenerit primo restituet omnia

et quomodo scriptum est in Filium hominis ut multa patia-

9:12 But I say unto you, That Elias is indeed come, and tur et contemnatur 9:12 sed dico uobis quia et Helias uenit

they have done unto him whatsoever they listed, as it is written of him.

et fecerunt illi quaecumque uoluerunt sicut scriptum est de

9:13 And when he came to his disciples, he saw a great eo 9:13 et ueniens ad discipulos suos uidit turbam magnam

multitude about them, and the scribes questioning with them.

circa eos et scribas conquirentes cum illis 9:14 et confes-

9:14 And straightway all the people, when they beheld him, were greatly amazed, and running to him saluted him.

tim omnis populus uidens eum stupefactus est et adcurren-

9:15 And he asked the scribes, What question ye with tes salutabant eum 9:15 et interrogauit eos quid inter uos

them?

9:16 And one of the multitude answered and said, Master, conquiritis 9:16 et respondens unus de turba dixit magister

I have brought unto thee my son, which hath a dumb spirit; 9:17 And wheresoever he taketh him, he teareth him: and adtuli filium meum ad te habentem spiritum mutum 9:17 he foameth, and gnasheth with his teeth, and pineth away: and I spake to thy disciples that they should cast him out;

and they could not. qui ubicumque eum adprehenderit adlidit eum et spumat et

stridet dentibus et arescit et dixi discipulis tuis ut eicerent

9:18 He answereth him, and saith, O faithless generation, illum et non potuerunt 9:18 qui respondens eis dicit o gehow long shall I be with you? how long shall I suffer you? bring him unto me.

neratio incredula quamdiu apud uos ero quamdiu uos patiar

9:19 And they brought him unto him: and when he saw adferte illum ad me 9:19 et adtulerunt eum et cum uidis-

him, straightway the spirit tare him; and he fell on the ground, and wallowed foaming.

set illum statim spiritus conturbauit eum et elisus in terram

9:20 And he asked his father, How long is it ago since this uolutabatur spumans 9:20 et interrogauit patrem eius quan-

came unto him? And he said, Of a child.

tum temporis est ex quo hoc ei accidit at ille ait ab infantia

9:21 And ofttimes it hath cast him into the fire, and into 9:21 et frequenter eum et in ignem et in aquas misit ut eum

the waters, to destroy him: but if thou canst do any thing, have compassion on us, and help us.

perderet sed si quid potes adiuua nos misertus nostri 9:22

9:22 Jesus said unto him, If thou canst believe, all things are possible to him that believeth.

Iesus autem ait illi si potes credere omnia possibilia credenti

9:23 And straightway the father of the child cried out, and 9:23 et continuo exclamans pater pueri cum lacrimis aiebat

said with tears, Lord, I believe; help thou mine unbelief.

9:24 When Jesus saw that the people came running to- credo adiuua incredulitatem meam 9:24 et cum uideret Ie-

gether, he rebuked the foul spirit, saying unto him, Thou dumb and deaf spirit, I charge thee, come out of him, and

enter no more into him. sus concurrentem turbam comminatus est spiritui inmundo

dicens illi surde et mute spiritus ego tibi praecipio exi ab

9:25 And the spirit cried, and rent him sore, and came out eo et amplius ne introeas in eum 9:25 et clamans et mulof him: and he was as one dead; insomuch that many said, He is dead.

Euangelium Secundum Marcum

tum discerpens eum exiit ab eo et factus est sicut mortuus ita ut multi dicerent quia mortuus est 9:26 Iesus autem te- 9:26 But Jesus took him by the hand, and lifted him up;

and he arose.

nens manum eius eleuauit illum et surrexit 9:27 et cum in- 9:27 And when he was come into the house, his disciples

asked him privately, Why could not we cast him out?

troisset in domum discipuli eius secreto interrogabant eum quare nos non potuimus eicere eum 9:28 et dixit illis hoc 9:28 And he said unto them, This kind can come forth by

nothing, but by prayer and fasting.

genus in nullo potest exire nisi in oratione et ieiunio 9:29 9:29 And they departed thence, and passed through

Galilee; and he would not that any man should know it.

et inde profecti praetergrediebantur Galilaeam nec uolebat quemquam scire 9:30 docebat autem discipulos suos et di- 9:30 For he taught his disciples, and said unto them, The

Son of man is delivered into the hands of men, and they shall kill him; and after that he is killed, he shall rise the

cebat illis quoniam Filius hominis tradetur in manus homi- third day. num et occident eum et occisus tertia die resurget 9:31 at 9:31 But they understood not that saying, and were afraid

to ask him.

illi ignorabant uerbum et timebant eum interrogare 9:32 et 9:32 And he came to Capernaum: and being in the house

he asked them, What was it that ye disputed among yourselves by the way?

uenerunt Capharnaum qui cum domi esset interrogabat eos quid in uia tractabatis 9:33 at illi tacebant siquidem inter se 9:33 But they held their peace: for by the way they had

disputed among themselves, who should be the greatest.

in uia disputauerant quis esset illorum maior 9:34 et resi- 9:34 And he sat down, and called the twelve, and saith

unto them, If any man desire to be first, the same shall be last of all, and servant of all.

dens uocauit duodecim et ait illis si quis uult primus esse erit omnium nouissimus et omnium minister 9:35 et accipiens 9:35 And he took a child, and set him in the midst of them:

and when he had taken him in his arms, he said unto them,

puerum statuit eum in medio eorum quem cum conplexus esset ait illis 9:36 quisquis unum ex huiusmodi pueris rece- 9:36 Whosoever shall receive one of such children in my

name, receiveth me: and whosoever shall receive me, receiveth not me, but him that sent me.

perit in nomine meo me recipit et quicumque me susceperit non me suscipit sed eum qui me misit 9:37 respondit illi Io- 9:37 And John answered him, saying, Master, we saw one

casting out devils in thy name, and he followeth not us: and we forbad him, because he followeth not us.

hannes dicens magister uidimus quendam in nomine tuo eicientem daemonia qui non sequitur nos et prohibuimus eum 9:38 Iesus autem ait nolite prohibere eum nemo est enim qui 9:38 But Jesus said, Forbid him not: for there is no man

which shall do a miracle in my name, that can lightly speak evil of me.

faciat uirtutem in nomine meo et possit cito male loqui de me 9:39 qui enim non est aduersum uos pro uobis est 9:40 9:39 For he that is not against us is on our part.

9:40 For whosoever shall give you a cup of water to drink in my name, because ye belong to Christ, verily I say unto you, he shall not lose his reward.

quisquis enim potum dederit uobis calicem aquae in nomine meo quia Christi estis amen dico uobis non perdet mercedem suam 9:41 et quisquis scandalizauerit unum ex his pusillis 9:41 And whosoever shall offend one of these little ones

that believe in me, it is better for him that a millstone were hanged about his neck, and he were cast into the sea.

credentibus in me bonum est ei magis si circumdaretur mola asinaria collo eius et in mare mitteretur 9:42 et si scanda- 9:42 And if thy hand offend thee, cut it off: it is better for

thee to enter into life maimed, than having two hands to go into hell, into the fire that never shall be quenched:

lizauerit te manus tua abscide illam bonum est tibi debilem introire in uitam quam duas manus habentem ire in gehennam in ignem inextinguibilem 9:43 ubi uermis eorum non 9:43 Where their worm dieth not, and the fire is not

quenched.

9:44 And if thy foot offend thee, cut it off: it is better for moritur et ignis non extinguitur 9:44 et si pes tuus te scan-

thee to enter halt into life, than having two feet to be cast into hell, into the fire that never shall be quenched:

dalizat amputa illum bonum est tibi claudum introire in uitam aeternam quam duos pedes habentem mitti in gehennam

9:45 Where their worm dieth not, and the fire is not ignis inextinguibilis 9:45 ubi uermis eorum non moritur et

quenched.

9:46 And if thine eye offend thee, pluck it out: it is better ignis non extinguitur 9:46 quod si oculus tuus scandalizat

for thee to enter into the kingdom of God with one eye, than having two eyes to be cast into hell fire:

te eice eum bonum est tibi luscum introire in regnum Dei

9:47 Where their worm dieth not, and the fire is not quam duos oculos habentem mitti in gehennam ignis 9:47

quenched.

9:48 For every one shall be salted with fire, and every sac- ubi uermis eorum non moritur et ignis non extinguitur 9:48

rifice shall be salted with salt.

9:49 Salt is good: but if the salt have lost his saltness, omnis enim igne sallietur et omnis uictima sallietur 9:49

wherewith will ye season it? Have salt in yourselves, and have peace one with another.

bonum est sal quod si sal insulsum fuerit in quo illud condietis habete in uobis sal et pacem habete inter uos

10:1 And he arose from thence, and cometh into the coasts 10:1 et inde exsurgens uenit in fines Iudaeae ultra Iorda-

of Judaea by the farther side of Jordan: and the people resort unto him again; and, as he was wont, he taught them

again. nen et conueniunt iterum turbae ad eum et sicut consueuerat 10:2 And the Pharisees came to him, and asked him, Is it iterum docebat illos 10:2 et accedentes Pharisaei interro-

lawful for a man to put away his wife? tempting him.

gabant eum si licet uiro uxorem dimittere temptantes eum

10:3 And he answered and said unto them, What did 10:3 at ille respondens dixit eis quid uobis praecepit Mo-

Moses command you?

10:4 And they said, Moses suffered to write a bill of di- ses 10:4 qui dixerunt Moses permisit libellum repudii sc-

vorcement, and to put her away.

10:5 And Jesus answered and said unto them, For the hard- ribere et dimittere 10:5 quibus respondens Iesus ait ad du-

ness of your heart he wrote you this precept.

10:6 But from the beginning of the creation God made ritiam cordis uestri scripsit uobis praeceptum istud 10:6 ab

them male and female.

initio autem creaturae masculum et feminam fecit eos Deus

10:7 For this cause shall a man leave his father and mother, 10:7 propter hoc relinquet homo patrem suum et matrem et and cleave to his wife; 10:8 And they twain shall be one flesh: so then they are no adherebit ad uxorem suam 10:8 et erunt duo in carne una more twain, but one flesh. 10:9 What therefore God hath joined together, let not man itaque iam non sunt duo sed una caro 10:9 quod ergo Deput asunder.

10:10 And in the house his disciples asked him again of us iunxit homo non separet 10:10 et in domo iterum disci-

the same matter.

10:11 And he saith unto them, Whosoever shall put away puli eius de eodem interrogauerunt eum 10:11 et dicit illis

his wife, and marry another, committeth adultery against her.

quicumque dimiserit uxorem suam et aliam duxerit adulte-

10:12 And if a woman shall put away her husband, and be rium committit super eam 10:12 et si uxor dimiserit uirum

married to another, she committeth adultery.

10:13 And they brought young children to him, that he suum et alii nupserit moechatur 10:13 et offerebant illi par-

should touch them: and his disciples rebuked those that brought them.

uulos ut tangeret illos discipuli autem comminabantur offe-

10:14 But when Jesus saw it, he was much displeased, and rentibus 10:14 quos cum uideret Iesus indigne tulit et ait

said unto them, Suffer the little children to come unto me, and forbid them not: for of such is the kingdom of God.

illis sinite paruulos uenire ad me et ne prohibueritis eos ta-

10:15 Verily I say unto you, Whosoever shall not receive lium est enim regnum Dei 10:15 amen dico uobis quisque the kingdom of God as a little child, he shall not enter therein.

non receperit regnum Dei uelut paruulus non intrabit in il-

Euangelium Secundum Marcum

lud 10:16 et conplexans eos et inponens manus super illos 10:16 And he took them up in his arms, put his hands upon

them, and blessed them.

benedicebat eos 10:17 et cum egressus esset in uiam pro- 10:17 And when he was gone forth into the way, there

came one running, and kneeled to him, and asked him, Good Master, what shall I do that I may inherit eternal

currens quidam genu flexo ante eum rogabat eum magister life? bone quid faciam ut uitam aeternam percipiam 10:18 Iesus 10:18 And Jesus said unto him, Why callest thou me

good? there is none good but one, that is, God.

autem dixit ei quid me dicis bonum nemo bonus nisi unus Deus 10:19 praecepta nosti ne adulteres ne occidas ne fu- 10:19 Thou knowest the commandments, Do not commit

adultery, Do not kill, Do not steal, Do not bear false witness, Defraud not, Honour thy father and mother.

reris ne falsum testimonium dixeris ne fraudem feceris honora patrem tuum et matrem 10:20 et ille respondens ait 10:20 And he answered and said unto him, Master, all

these have I observed from my youth.

illi magister omnia haec conseruaui a iuuentute mea 10:21 10:21 Then Jesus beholding him loved him, and said unto

him, One thing thou lackest: go thy way, sell whatsoever thou hast, and give to the poor, and thou shalt have treasure

Iesus autem intuitus eum dilexit eum et dixit illi unum tibi in heaven: and come, take up the cross, and follow me. deest uade quaecumque habes uende et da pauperibus et habebis thesaurum in caelo et ueni sequere me 10:22 qui con- 10:22 And he was sad at that saying, and went away

grieved: for he had great possessions.

tristatus in uerbo abiit maerens erat enim habens possessiones multas 10:23 et circumspiciens Iesus ait discipulis suis 10:23 And Jesus looked round about, and saith unto his

disciples, How hardly shall they that have riches enter into the kingdom of God!

quam difficile qui pecunias habent in regnum Dei introibunt 10:24 discipuli autem obstupescebant in uerbis eius at Iesus 10:24 And the disciples were astonished at his words. But

Jesus answereth again, and saith unto them, Children, how hard is it for them that trust in riches to enter into the king-

rursus respondens ait illis filioli quam difficile est confiden- dom of God! tes in pecuniis regnum Dei introire 10:25 facilius est ca- 10:25 It is easier for a camel to go through the eye of a

needle, than for a rich man to enter into the kingdom of God.

melum per foramen acus transire quam diuitem intrare in regnum Dei 10:26 qui magis admirabantur dicentes ad se- 10:26 And they were astonished out of measure, saying

among themselves, Who then can be saved?

met ipsos et quis potest saluus fieri 10:27 et intuens illos 10:27 And Jesus looking upon them saith, With men it is

impossible, but not with God: for with God all things are possible.

Iesus ait apud homines inpossibile est sed non apud Deum omnia enim possibilia sunt apud Deum 10:28 coepit Pet- 10:28 Then Peter began to say unto him, Lo, we have left

all, and have followed thee.

rus ei dicere ecce nos dimisimus omnia et secuti sumus te 10:29 respondens Iesus ait amen dico uobis nemo est qui re- 10:29 And Jesus answered and said, Verily I say unto you,

There is no man that hath left house, or brethren, or sisters, or father, or mother, or wife, or children, or lands, for my

liquerit domum aut fratres aut sorores aut matrem aut patrem sake, and the gospel’s, aut filios aut agros propter me et propter euangelium 10:30 10:30 But he shall receive an hundredfold now in this time,

houses, and brethren, and sisters, and mothers, and children, and lands, with persecutions; and in the world to

qui non accipiat centies tantum nunc in tempore hoc domos come eternal life. et fratres et sorores et matres et filios et agros cum persecutionibus et in saeculo futuro uitam aeternam 10:31 multi 10:31 But many that are first shall be last; and the last first. autem erunt primi nouissimi et nouissimi primi 10:32 erant 10:32 And they were in the way going up to Jerusalem;

and Jesus went before them: and they were amazed; and as they followed, they were afraid. And he took again the twelve, and began to tell them what things should happen unto him,

autem in uia ascendentes in Hierosolyma et praecedebat illos Iesus et stupebant et sequentes timebant et adsumens iterum

Biblia Sacra Vulgata

10:33 Saying, Behold, we go up to Jerusalem; and the Son duodecim coepit illis dicere quae essent ei uentura 10:33

of man shall be delivered unto the chief priests, and unto the scribes; and they shall condemn him to death, and shall

deliver him to the Gentiles: quia ecce ascendimus in Hierosolyma et Filius hominis tra-

detur principibus sacerdotum et scribis et senioribus et dam-

10:34 And they shall mock him, and shall scourge him, nabunt eum morti et tradent eum gentibus 10:34 et inludent

and shall spit upon him, and shall kill him: and the third day he shall rise again.

ei et conspuent eum et flagellabunt eum et interficient eum

10:35 And James and John, the sons of Zebedee, come et tertia die resurget 10:35 et accedunt ad illum Iacobus et

unto him, saying, Master, we would that thou shouldest do for us whatsoever we shall desire.

Iohannes filii Zebedaei dicentes magister uolumus ut quod-

10:36 And he said unto them, What would ye that I should cumque petierimus facias nobis 10:36 at ille dixit eis quid

do for you?

10:37 They said unto him, Grant unto us that we may sit, uultis ut faciam uobis 10:37 et dixerunt da nobis ut unus

one on thy right hand, and the other on thy left hand, in thy glory.

ad dexteram tuam et alius ad sinistram tuam sedeamus in

10:38 But Jesus said unto them, Ye know not what ye ask: gloria tua 10:38 Iesus autem ait eis nescitis quid petatis po-

can ye drink of the cup that I drink of? and be baptized with the baptism that I am baptized with?

testis bibere calicem quem ego bibo aut baptismum quo ego

10:39 And they said unto him, We can. And Jesus said baptizor baptizari 10:39 at illi dixerunt ei possumus Iesus

unto them, Ye shall indeed drink of the cup that I drink of; and with the baptism that I am baptized withal shall ye be

baptized: autem ait eis calicem quidem quem ego bibo bibetis et bap-10:40 But to sit on my right hand and on my left hand is tismum quo ego baptizor baptizabimini 10:40 sedere autem

not mine to give; but it shall be given to them for whom it is prepared.

ad dexteram meam uel ad sinistram non est meum dare sed

10:41 And when the ten heard it, they began to be much quibus paratum est 10:41 et audientes decem coeperunt in-

displeased with James and John.

10:42 But Jesus called them to him, and saith unto them, dignari de Iacobo et Iohanne 10:42 Iesus autem uocans eos

Ye know that they which are accounted to rule over the Gentiles exercise lordship over them; and their great ones

exercise authority upon them. ait illis scitis quia hii qui uidentur principari gentibus do-

minantur eis et principes eorum potestatem habent ipsorum

10:43 But so shall it not be among you: but whosoever 10:43 non ita est autem in uobis sed quicumque uoluerit fieri

will be great among you, shall be your minister:

10:44 And whosoever of you will be the chiefest, shall be maior erit uester minister 10:44 et quicumque uoluerit in

servant of all.

10:45 For even the Son of man came not to be minis- uobis primus esse erit omnium seruus 10:45 nam et Filius

tered unto, but to minister, and to give his life a ransom for many.

hominis non uenit ut ministraretur ei sed ut ministraret et

10:46 And they came to Jericho: and as he went out of daret animam suam redeeptionem pro multis 10:46 et ue-

Jericho with his disciples and a great number of people, blind Bartimaeus, the son of Timaeus, sat by the highway

side begging. niunt Hierichum et proficiscente eo de Hiericho et discipulis

eius et plurima multitudine filius Timei Bartimeus caecus

10:47 And when he heard that it was Jesus of Nazareth, he sedebat iuxta uiam mendicans 10:47 qui cum audisset quia

began to cry out, and say, Jesus, thou son of David, have mercy on me.

Iesus Nazarenus est coepit clamare et dicere Fili Dauid Iesu

10:48 And many charged him that he should hold his miserere mei 10:48 et comminabantur illi multi ut taceret at

peace: but he cried the more a great deal, Thou son of David, have mercy on me.

ille multo magis clamabat Fili Dauid miserere mei 10:49 et

10:49 And Jesus stood still, and commanded him to be called. And they call the blind man, saying unto him, Be of good comfort, rise; he calleth thee.

stans Iesus praecepit illum uocari et uocant caecum dicentes

10:50 And he, casting away his garment, rose, and came ei animaequior esto surge uocat te 10:50 qui proiecto ues-

to Jesus.

Euangelium Secundum Marcum

timento suo exiliens uenit ad eum 10:51 et respondens illi 10:51 And Jesus answered and said unto him, What wilt

thou that I should do unto thee? The blind man said unto him, Lord, that I might receive my sight.

Iesus dixit quid uis tibi faciam caecus autem dixit ei rabboni ut uideam 10:52 Iesus autem ait illi uade fides tua te saluum 10:52 And Jesus said unto him, Go thy way; thy faith hath

made thee whole. And immediately he received his sight, and followed Jesus in the way.

fecit et confestim uidit et sequebatur eum in uia

11:1 et cum adpropinquarent Hierosolymae et Bethaniae 11:1 And when they came nigh to Jerusalem, unto Beth-

phage and Bethany, at the mount of Olives, he sendeth forth two of his disciples,

ad montem Oliuarum mittit duos ex discipulis suis 11:2 et

11:2 And saith unto them, Go your way into the village over against you: and as soon as ye be entered into it, ye shall find a colt tied, whereon never man sat; loose him, and bring him.

ait illis ite in castellum quod est contra uos et statim introeuntes illuc inuenietis pullum ligatum super quem nemo adhuc hominum sedit soluite illum et adducite 11:3 et si 11:3 And if any man say unto you, Why do ye this? say

ye that the Lord hath need of him; and straightway he will send him hither.

quis uobis dixerit quid facitis dicite quia Domino necessarius est et continuo illum dimittet huc 11:4 et abeuntes inuene- 11:4 And they went their way, and found the colt tied by

the door without in a place where two ways met; and they loose him.

runt pullum ligatum ante ianuam foris in biuio et soluunt eum 11:5 et quidam de illic stantibus dicebant illis quid fa- 11:5 And certain of them that stood there said unto them,

What do ye, loosing the colt?

citis soluentes pullum 11:6 qui dixerunt eis sicut praece- 11:6 And they said unto them even as Jesus had com-

manded: and they let them go.

perat illis Iesus et dimiserunt eis 11:7 et duxerunt pullum 11:7 And they brought the colt to Jesus, and cast their gar-

ments on him; and he sat upon him.

ad Iesum et inponunt illi uestimenta sua et sedit super eo 11:8 multi autem uestimenta sua strauerunt in uia alii autem 11:8 And many spread their garments in the way: and oth-

ers cut down branches off the trees, and strawed them in the way.

frondes caedebant de arboribus et sternebant in uia 11:9 et

11:9 And they that went before, and they that followed, cried, saying, Hosanna; Blessed is he that cometh in the name of the Lord:

qui praeibant et qui sequebantur clamabant dicentes osanna benedictus qui uenit in nomine Domini 11:10 benedictum 11:10 Blessed be the kingdom of our father David, that

cometh in the name of the Lord: Hosanna in the highest.

quod uenit regnum patris nostri Dauid osanna in excelsis 11:11 et introiuit Hierosolyma in templum et circumspectis 11:11 And Jesus entered into Jerusalem, and into the tem-

ple: and when he had looked round about upon all things, and now the eventide was come, he went out unto Bethany

omnibus cum iam uespera esset hora exiuit in Bethania cum with the twelve. duodecim 11:12 et alia die cum exirent a Bethania esuriit 11:12 And on the morrow, when they were come from

Bethany, he was hungry:

11:13 cumque uidisset a longe ficum habentem folia uenit si 11:13 And seeing a fig tree afar off having leaves, he came,

if haply he might find any thing thereon: and when he came to it, he found nothing but leaves; for the time of figs

quid forte inueniret in ea et cum uenisset ad eam nihil in- was not yet. uenit praeter folia non enim erat tempus ficorum 11:14 et 11:14 And Jesus answered and said unto it, No man eat

fruit of thee hereafter for ever. And his disciples heard it.

respondens dixit ei iam non amplius in aeternum quisquam fructum ex te manducet et audiebant discipuli eius 11:15 11:15 And they come to Jerusalem: and Jesus went into

the temple, and began to cast out them that sold and bought in the temple, and overthrew the tables of the money-

et ueniunt Hierosolymam et cum introisset templum coepit changers, and the seats of them that sold doves; eicere uendentes et ementes in templo et mensas nummulariorum et cathedras uendentium columbas euertit 11:16 et 11:16 And would not suffer that any man should carry any

vessel through the temple.

non sinebat ut quisquam uas transferret per templum 11:17 11:17 And he taught, saying unto them, Is it not writ-

ten, My house shall be called of all nations the house of prayer? but ye have made it a den of thieves.

Biblia Sacra Vulgata

et docebat dicens eis non scriptum est quia domus mea domus orationis uocabitur omnibus gentibus uos autem fecistis

11:18 And the scribes and chief priests heard it, and sought eam speluncam latronum 11:18 quo audito principes sacerhow they might destroy him: for they feared him, because all the people was astonished at his doctrine.

dotum et scribae quaerebant quomodo eum perderent timebant enim eum quoniam uniuersa turba admirabatur super

11:19 And when even was come, he went out of the city. doctrina eius 11:19 et cum uespera facta esset egredieba-11:20 And in the morning, as they passed by, they saw the tur de ciuitate 11:20 et cum mane transirent uiderunt ficum

fig tree dried up from the roots.

11:21 And Peter calling to remembrance saith unto him, aridam factam a radicibus 11:21 et recordatus Petrus dicit

Master, behold, the fig tree which thou cursedst is withered away.

ei rabbi ecce ficus cui maledixisti aruit 11:22 et respondens

11:22 And Jesus answering saith unto them, Have faith in God.

11:23 For verily I say unto you, That whosoever shall say Iesus ait illis habete fidem Dei 11:23 amen dico uobis qui-

unto this mountain, Be thou removed, and be thou cast into the sea; and shall not doubt in his heart, but shall believe that those things which he saith shall come to pass; he shall have whatsoever he saith.

cumque dixerit huic monti tollere et mittere in mare et non haesitauerit in corde suo sed crediderit quia quodcumque

11:24 Therefore I say unto you, What things soever ye dixerit fiat fiet ei 11:24 propterea dico uobis omnia quae-

desire, when ye pray, believe that ye receive them, and ye shall have them.

cumque orantes petitis credite quia accipietis et ueniet uobis

11:25 And when ye stand praying, forgive, if ye have 11:25 et cum stabitis ad orandum dimittite si quid habetis ad-

ought against any: that your Father also which is in heaven may forgive you your trespasses.

uersus aliquem ut et Pater uester qui in caelis est dimittat

11:26 But if ye do not forgive, neither will your Father uobis peccata uestra 11:26 quod si uos non dimiseritis nec

which is in heaven forgive your trespasses.

Pater uester qui in caelis est dimittet uobis peccata uestra

11:27 And they come again to Jerusalem: and as he was 11:27 et ueniunt rursus Hierosolymam et cum ambularet in

walking in the temple, there come to him the chief priests, and the scribes, and the elders,

templo accedunt ad eum summi sacerdotes et scribae et se-

11:28 And say unto him, By what authority doest thou niores 11:28 et dicunt illi in qua potestate haec facis et quis

these things? and who gave thee this authority to do these things?

tibi dedit hanc potestatem ut ista facias 11:29 Iesus autem

11:29 And Jesus answered and said unto them, I will also ask of you one question, and answer me, and I will tell you by what authority I do these things.

respondens ait illis interrogabo uos et ego unum uerbum et respondete mihi et dicam uobis in qua potestate haec faciam

11:30 The baptism of John, was it from heaven, or of men? 11:30 baptismum Iohannis de caelo erat an ex hominibus res-

answer me.

11:31 And they reasoned with themselves, saying, If we pondete mihi 11:31 at illi cogitabant secum dicentes si dixe-

shall say, From heaven; he will say, Why then did ye not believe him?

rimus de caelo dicet quare ergo non credidistis ei 11:32 sed

11:32 But if we shall say, Of men; they feared the people: for all men counted John, that he was a prophet indeed.

dicemus ex hominibus timebant populum omnes enim ha-

11:33 And they answered and said unto Jesus, We cannot bebant Iohannem quia uere propheta esset 11:33 et respon-

tell. And Jesus answering saith unto them, Neither do I tell you by what authority I do these things.

dentes dicunt Iesu nescimus respondens Iesus ait illis neque ego dico uobis in qua potestate haec faciam

12:1 And he began to speak unto them by parables. A 12:1 et coepit illis in parabolis loqui uineam pastinauit

certain man planted a vineyard, and set an hedge about it, and digged a place for the winefat, and built a tower, and

let it out to husbandmen, and went into a far country. homo et circumdedit sepem et fodit lacum et aedificauit tur-

Euangelium Secundum Marcum

rem et locauit eam agricolis et peregre profectus est 12:2 12:2 And at the season he sent to the husbandmen a ser-

vant, that he might receive from the husbandmen of the fruit of the vineyard.

et misit ad agricolas in tempore seruum ut ab agricolis acciperet de fructu uineae 12:3 qui adprehensum eum ceci- 12:3 And they caught him, and beat him, and sent him

away empty.

derunt et dimiserunt uacuum 12:4 et iterum misit ad illos 12:4 And again he sent unto them another servant; and at

him they cast stones, and wounded him in the head, and sent him away shamefully handled.

alium seruum et illum capite uulnerauerunt et contumeliis adfecerunt 12:5 et rursum alium misit et illum occiderunt et 12:5 And again he sent another; and him they killed, and

many others; beating some, and killing some.

plures alios quosdam caedentes alios uero occidentes 12:6 12:6 Having yet therefore one son, his wellbeloved, he

sent him also last unto them, saying, They will reverence my son.

adhuc ergo unum habens filium carissimum et illum misit ad eos nouissimum dicens quia reuerebuntur filium meum 12:7 coloni autem dixerunt ad inuicem hic est heres uenite 12:7 But those husbandmen said among themselves, This

is the heir; come, let us kill him, and the inheritance shall be our’s.

occidamus eum et nostra erit hereditas 12:8 et adprehen-

12:8 And they took him, and killed him, and cast him out of the vineyard.

dentes eum occiderunt et eiecerunt extra uineam 12:9 quid 12:9 What shall therefore the lord of the vineyard do? he

will come and destroy the husbandmen, and will give the vineyard unto others.

ergo faciet dominus uineae ueniet et perdet colonos et dabit uineam aliis 12:10 nec scripturam hanc legistis lapidem 12:10 And have ye not read this scripture; The stone which

the builders rejected is become the head of the corner:

quem reprobauerunt aedificantes hic factus est in caput an-

guli 12:11 a Domino factum est istud et est mirabile in 12:11 This was the Lord’s doing, and it is marvellous in

our eyes?

oculis nostris 12:12 et quaerebant eum tenere et timuerunt 12:12 And they sought to lay hold on him, but feared

the people: for they knew that he had spoken the parable against them: and they left him, and went their way.

turbam cognouerunt enim quoniam ad eos parabolam hanc dixerit et relicto eo abierunt 12:13 et mittunt ad eum quos- 12:13 And they send unto him certain of the Pharisees and

of the Herodians, to catch him in his words.

dam ex Pharisaeis et Herodianis ut eum caperent in uerbo 12:14 qui uenientes dicunt ei magister scimus quoniam uerax 12:14 And when they were come, they say unto him, Mas-

ter, we know that thou art true, and carest for no man: for thou regardest not the person of men, but teachest the way of God in truth: Is it lawful to give tribute to Caesar, or not?

es et non curas quemquam nec enim uides in faciem hominis sed in ueritate uiam Dei doces licet dari tributum Caesari an non dabimus 12:15 qui sciens uersutiam eorum ait illis quid 12:15 Shall we give, or shall we not give? But he, knowing

their hypocrisy, said unto them, Why tempt ye me? bring me a penny, that I may see it.

me temptatis adferte mihi denarium ut uideam 12:16 at illi

12:16 And they brought it. And he saith unto them, Whose is this image and superscription? And they said unto him, Caesar’s.

adtulerunt et ait illis cuius est imago haec et inscriptio dicunt illi Caesaris 12:17 respondens autem Iesus dixit illis reddite 12:17 And Jesus answering said unto them, Render to

Caesar the things that are Caesar’s, and to God the things that are God’s. And they marvelled at him.

igitur quae sunt Caesaris Caesari et quae sunt Dei Deo et mirabantur super eo 12:18 et uenerunt ad eum Sadducaei qui 12:18 Then come unto him the Sadducees, which say there

is no resurrection; and they asked him, saying,

dicunt resurrectionem non esse et interrogabant eum dicentes 12:19 magister Moses nobis scripsit ut si cuius frater 12:19 Master, Moses wrote unto us, If a man’s brother die,

and leave his wife behind him, and leave no children, that his brother should take his wife, and raise up seed unto his

mortuus fuerit et dimiserit uxorem et filios non reliquerit ac- brother. cipiat frater eius uxorem ipsius et resuscitet semen fratri suo

Biblia Sacra Vulgata

12:20 Now there were seven brethren: and the first took a 12:20 septem ergo fratres erant et primus accepit uxorem et

wife, and dying left no seed.

12:21 And the second took her, and died, neither left he mortuus est non relicto semine 12:21 et secundus accepit

any seed: and the third likewise.

eam et mortuus est et nec iste reliquit semen et tertius si-

12:22 And the seven had her, and left no seed: last of all militer 12:22 et acceperunt eam similiter septem et non re-

the woman died also.

liquerunt semen nouissima omnium defuncta est et mulier

12:23 In the resurrection therefore, when they shall rise, 12:23 in resurrectione ergo cum resurrexerint cuius de his erit

whose wife shall she be of them? for the seven had her to wife.

uxor septem enim habuerunt eam uxorem 12:24 et respon-

12:24 And Jesus answering said unto them, Do ye not therefore err, because ye know not the scriptures, neither the power of God?

dens Iesus ait illis non ideo erratis non scientes scripturas

12:25 For when they shall rise from the dead, they nei- neque uirtutem Dei 12:25 cum enim a mortuis resurrexerint

ther marry, nor are given in marriage; but are as the angels which are in heaven.

neque nubent neque nubentur sed sunt sicut angeli in caelis

12:26 And as touching the dead, that they rise: have ye 12:26 de mortuis autem quod resurgant non legistis in libro

not read in the book of Moses, how in the bush God spake unto him, saying, I am the God of Abraham, and the God

of Isaac, and the God of Jacob? Mosi super rubum quomodo dixerit illi Deus inquiens ego 12:27 He is not the God of the dead, but the God of the sum Deus Abraham et Deus Isaac et Deus Iacob 12:27 non

living: ye therefore do greatly err.

est Deus mortuorum sed uiuorum uos ergo multum erratis

12:28 And one of the scribes came, and having heard them 12:28 et accessit unus de scribis qui audierat illos conquiren-

reasoning together, and perceiving that he had answered them well, asked him, Which is the first commandment of

all? tes et uidens quoniam bene illis responderit interrogauit eum

12:29 And Jesus answered him, The first of all the com- quod esset primum omnium mandatum 12:29 Iesus autem mandments is, Hear, O Israel; The Lord our God is one Lord:

respondit ei quia primum omnium mandatum est audi Is-

12:30 And thou shalt love the Lord thy God with all thy rahel Dominus Deus noster Deus unus est 12:30 et diliges

heart, and with all thy soul, and with all thy mind, and with all thy strength: this is the first commandment.

Dominum Deum tuum ex toto corde tuo et ex tota anima tua et ex tota mente tua et ex tota uirtute tua hoc est pri-

12:31 And the second is like, namely this, Thou shalt love mum mandatum 12:31 secundum autem simile illi diliges thy neighbour as thyself. There is none other commandment greater than these.

proximum tuum tamquam te ipsum maius horum aliud man-

12:32 And the scribe said unto him, Well, Master, thou datum non est 12:32 et ait illi scriba bene magister in ueri-

hast said the truth: for there is one God; and there is none other but he:

tate dixisti quia unus est et non est alius praeter eum 12:33

12:33 And to love him with all the heart, and with all the understanding, and with all the soul, and with all the strength, and to love his neighbour as himself, is more than all whole burnt offerings and sacrifices.

et ut diligatur ex toto corde et ex toto intellectu et ex tota anima et ex tota fortitudine et diligere proximum tamquam se ipsum maius est omnibus holocaustomatibus et sacrificiis

12:34 And when Jesus saw that he answered discreetly, he 12:34 Iesus autem uidens quod sapienter respondisset dixit

said unto him, Thou art not far from the kingdom of God. And no man after that durst ask him any question.

illi non es longe a regno Dei et nemo iam audebat eum inter-

12:35 And Jesus answered and said, while he taught in rogare 12:35 et respondens Iesus dicebat docens in templo

the temple, How say the scribes that Christ is the son of David?

quomodo dicunt scribae Christum Filium esse Dauid 12:36

12:36 For David himself said by the Holy Ghost, The LORD said to my Lord, Sit thou on my right hand, till I make thine enemies thy footstool.

ipse enim Dauid dicit in Spiritu Sancto dixit Dominus Do-

Euangelium Secundum Marcum

mino meo sede a dextris meis donec ponam inimicos tuos scabillum pedum tuorum 12:37 ipse ergo Dauid dicit eum 12:37 David therefore himself calleth him Lord; and

whence is he then his son? And the common people heard him gladly.

Dominum et unde est filius eius et multa turba eum libenter audiuit 12:38 et dicebat eis in doctrina sua cauete a scri- 12:38 And he said unto them in his doctrine, Beware of

the scribes, which love to go in long clothing, and love salutations in the marketplaces,

bis qui uolunt in stolis ambulare et salutari in foro 12:39 et

12:39 And the chief seats in the synagogues, and the uppermost rooms at feasts:

in primis cathedris sedere in synagogis et primos discubitus in cenis 12:40 qui deuorant domos uiduarum sub obtentu 12:40 Which devour widows’ houses, and for a pretence

make long prayers: these shall receive greater damnation.

prolixae orationis hii accipient prolixius iudicium 12:41 et 12:41 And Jesus sat over against the treasury, and beheld

how the people cast money into the treasury: and many that were rich cast in much.

sedens Iesus contra gazofilacium aspiciebat quomodo turba iactaret aes in gazofilacium et multi diuites iactabant multa 12:42 cum uenisset autem una uidua pauper misit duo minuta 12:42 And there came a certain poor widow, and she threw

in two mites, which make a farthing.

quod est quadrans 12:43 et conuocans discipulos suos ait 12:43 And he called unto him his disciples, and saith unto

them, Verily I say unto you, That this poor widow hath cast more in, than all they which have cast into the treasury:

illis amen dico uobis quoniam uidua haec pauper plus omnibus misit qui miserunt in gazofilacium 12:44 omnes enim 12:44 For all they did cast in of their abundance; but she

of her want did cast in all that she had, even all her living.

ex eo quod abundabat illis miserunt haec uero de penuria sua omnia quae habuit misit totum uictum suum

13:1 et cum egrederetur de templo ait illi unus ex disci- 13:1 And as he went out of the temple, one of his disciples

saith unto him, Master, see what manner of stones and what buildings are here!

pulis suis magister aspice quales lapides et quales structurae 13:2 et respondens Iesus ait illi uides has omnes magnas 13:2 And Jesus answering said unto him, Seest thou these

great buildings? there shall not be left one stone upon another, that shall not be thrown down.

aedificationes non relinquetur lapis super lapidem qui non destruatur 13:3 et cum sederet in montem Oliuarum con- 13:3 And as he sat upon the mount of Olives over against

the temple, Peter and James and John and Andrew asked him privately,

tra templum interrogabant eum separatim Petrus et Iacobus et Iohannes et Andreas 13:4 dic nobis quando ista fient et 13:4 Tell us, when shall these things be? and what shall

be the sign when all these things shall be fulfilled?

quod signum erit quando haec omnia incipient consummari 13:5 et respondens Iesus coepit dicere illis uidete ne quis uos 13:5 And Jesus answering them began to say, Take heed

lest any man deceive you:

seducat 13:6 multi enim uenient in nomine meo dicentes 13:6 For many shall come in my name, saying, I am

Christ; and shall deceive many.

quia ego sum et multos seducent 13:7 cum audieritis au- 13:7 And when ye shall hear of wars and rumours of wars,

be ye not troubled: for such things must needs be; but the end shall not be yet.

tem bella et opiniones bellorum ne timueritis oportet enim fieri sed nondum finis 13:8 exsurget autem gens super gen- 13:8 For nation shall rise against nation, and kingdom

against kingdom: and there shall be earthquakes in divers places, and there shall be famines and troubles: these are

tem et regnum super regnum et erunt terraemotus per loca the beginnings of sorrows.

et fames initium dolorum haec 13:9 uidete autem uosmet 13:9 But take heed to yourselves: for they shall deliver you up to councils; and in the synagogues ye shall be beaten: and ye shall be brought before rulers and kings for my

ipsos tradent enim uos conciliis et in synagogis uapulabi- sake, for a testimony against them. tis et ante praesides et reges stabitis propter me in testimo-

Biblia Sacra Vulgata

13:10 And the gospel must first be published among all nium illis 13:10 et in omnes gentes primum oportet praedi-

nations.

13:11 But when they shall lead you, and deliver you up, cari euangelium 13:11 et cum duxerint uos tradentes nolite

take no thought beforehand what ye shall speak, neither do ye premeditate: but whatsoever shall be given you in that hour, that speak ye: for it is not ye that speak, but the Holy Ghost.

praecogitare quid loquamini sed quod datum uobis fuerit in illa hora id loquimini non enim estis uos loquentes sed Spi-

13:12 Now the brother shall betray the brother to death, ritus Sanctus 13:12 tradet autem frater fratrem in mortem et

and the father the son; and children shall rise up against their parents, and shall cause them to be put to death.

pater filium et consurgent filii in parentes et morte adficient

13:13 And ye shall be hated of all men for my name’s eos 13:13 et eritis odio omnibus propter nomen meum qui

sake: but he that shall endure unto the end, the same shall be saved.

autem sustinuerit in finem hic saluus erit 13:14 cum autem

13:14 But when ye shall see the abomination of desolation, spoken of by Daniel the prophet, standing where it ought not, (let him that readeth understand,) then let them that be in Judaea flee to the mountains:

uideritis abominationem desolationis stantem ubi non debet qui legit intellegat tunc qui in Iudaea sunt fugiant in montes

13:15 And let him that is on the housetop not go down into 13:15 et qui super tectum ne descendat in domum nec introeat

the house, neither enter therein, to take any thing out of his house:

ut tollat quid de domo sua 13:16 et qui in agro erit non re-

13:16 And let him that is in the field not turn back again for to take up his garment.

13:17 But woe to them that are with child, and to them that uertatur retro tollere uestimentum suum 13:17 uae autem

give suck in those days!

13:18 And pray ye that your flight be not in the winter. praegnatibus et nutrientibus in illis diebus 13:18 orate uero 13:19 For in those days shall be affliction, such as was not ut hieme non fiant 13:19 erunt enim dies illi tribulationes

from the beginning of the creation which God created unto this time, neither shall be.

tales quales non fuerunt ab initio creaturae quam condidit

13:20 And except that the Lord had shortened those days, Deus usque nunc neque fient 13:20 et nisi breuiasset Dono flesh should be saved: but for the elect’s sake, whom he hath chosen, he hath shortened the days.

minus dies non fuisset salua omnis caro sed propter electos

13:21 And then if any man shall say to you, Lo, here is quos elegit breuiauit dies 13:21 et tunc si quis uobis dixerit

Christ; or, lo, he is there; believe him not:

13:22 For false Christs and false prophets shall rise, and ecce hic est Christus ecce illic ne credideritis 13:22 exsur-

shall shew signs and wonders, to seduce, if it were possible, even the elect.

gent enim pseudochristi et pseudoprophetae et dabunt signa

13:23 But take ye heed: behold, I have foretold you all et portenta ad seducendos si potest fieri etiam electos 13:23

things. 13:24 But in those days, after that tribulation, the sun shall uos ergo uidete ecce praedixi uobis omnia 13:24 sed in illis be darkened, and the moon shall not give her light,

diebus post tribulationem illam sol contenebrabitur et luna

13:25 And the stars of heaven shall fall, and the powers non dabit splendorem suum 13:25 et erunt stellae caeli de-

that are in heaven shall be shaken.

13:26 And then shall they see the Son of man coming in cidentes et uirtutes quae sunt in caelis mouebuntur 13:26 et

the clouds with great power and glory.

tunc uidebunt Filium hominis uenientem in nubibus cum uir-

13:27 And then shall he send his angels, and shall gather tute multa et gloria 13:27 et tunc mittet angelos suos et con-

together his elect from the four winds, from the uttermost part of the earth to the uttermost part of heaven.

gregabit electos suos a quattuor uentis a summo terrae usque

13:28 Now learn a parable of the fig tree; When her branch ad summum caeli 13:28 a ficu autem discite parabolam cum is yet tender, and putteth forth leaves, ye know that summer is near:

iam ramus eius tener fuerit et nata fuerint folia cognosci-

13:29 So ye in like manner, when ye shall see these things tis quia in proximo sit aestas 13:29 sic et uos cum uideritis

come to pass, know that it is nigh, even at the doors. 13:30 Verily I say unto you, that this generation shall not haec fieri scitote quod in proximo sit in ostiis 13:30 amen pass, till all these things be done.

Euangelium Secundum Marcum

dico uobis quoniam non transiet generatio haec donec omnia ista fiant 13:31 caelum et terra transibunt uerba autem mea 13:31 Heaven and earth shall pass away: but my words

shall not pass away.

non transibunt 13:32 de die autem illo uel hora nemo scit 13:32 But of that day and that hour knoweth no man, no,

not the angels which are in heaven, neither the Son, but the Father.

neque angeli in caelo neque Filius nisi Pater 13:33 uidete

13:33 Take ye heed, watch and pray: for ye know not when the time is.

uigilate et orate nescitis enim quando tempus sit 13:34 sic- 13:34 For the Son of Man is as a man taking a far jour-

ney, who left his house, and gave authority to his servants, and to every man his work, and commanded the porter to

ut homo qui peregre profectus reliquit domum suam et dedit watch. seruis suis potestatem cuiusque operis et ianitori praecipiat ut uigilet 13:35 uigilate ergo nescitis enim quando dominus 13:35 Watch ye therefore: for ye know not when the mas-

ter of the house cometh, at even, or at midnight, or at the cockcrowing, or in the morning:

domus ueniat sero an media nocte an galli cantu an mane 13:36 ne cum uenerit repente inueniat uos dormientes 13:37 13:36 Lest coming suddenly he find you sleeping.

13:37 And what I say unto you I say unto all, Watch.

quod autem uobis dico omnibus dico uigilate

14:1 erat autem pascha et azyma post biduum et quaere- 14:1 After two days was the feast of the passover, and of unleavened bread: and the chief priests and the scribes sought how they might take him by craft, and put him to

bant summi sacerdotes et scribae quomodo eum dolo tene- death. rent et occiderent 14:2 dicebant enim non in die festo ne 14:2 But they said, Not on the feast day, lest there be an

uproar of the people.

forte tumultus fieret populi 14:3 et cum esset Bethaniae in 14:3 And being in Bethany in the house of Simon the leper,

as he sat at meat, there came a woman having an alabaster box of ointment of spikenard very precious; and she brake

domo Simonis leprosi et recumberet uenit mulier habens ala- the box, and poured it on his head. bastrum unguenti nardi spicati pretiosi et fracto alabastro effudit super caput eius 14:4 erant autem quidam indigne fe- 14:4 And there were some that had indignation within

themselves, and said, Why was this waste of the ointment made?

rentes intra semet ipsos et dicentes ut quid perditio ista unguenti facta est 14:5 poterat enim unguentum istud ueniri 14:5 For it might have been sold for more than three hun-

dred pence, and have been given to the poor. And they murmured against her.

plus quam trecentis denariis et dari pauperibus et fremebant in eam 14:6 Iesus autem dixit sinite eam quid illi molesti 14:6 And Jesus said, Let her alone; why trouble ye her?

she hath wrought a good work on me.

estis bonum opus operata est in me 14:7 semper enim pau- 14:7 For ye have the poor with you always, and whenso-

ever ye will ye may do them good: but me ye have not always.

peres habetis uobiscum et cum uolueritis potestis illis benefacere me autem non semper habetis 14:8 quod habuit haec 14:8 She hath done what she could: she is come aforehand

to anoint my body to the burying.

fecit praeuenit unguere corpus meum in sepulturam 14:9 14:9 Verily I say unto you, Wheresoever this gospel shall be preached throughout the whole world, this also that she hath done shall be spoken of for a memorial of her.

amen dico uobis ubicumque praedicatum fuerit euangelium istud in uniuersum mundum et quod fecit haec narrabitur in

memoriam eius 14:10 et Iudas Scariotis unus de duodecim 14:10 And Judas Iscariot, one of the twelve, went unto the

chief priests, to betray him unto them.

abiit ad summos sacerdotes ut prodeeet eum illis 14:11 qui 14:11 And when they heard it, they were glad, and

promised to give him money. And he sought how he might conveniently betray him.

audientes gauisi sunt et promiserunt ei pecuniam se daturos et quaerebat quomodo illum oportune traderet 14:12 et 14:12 And the first day of unleavened bread, when they

killed the passover, his disciples said unto him, Where wilt thou that we go and prepare that thou mayest eat the

primo die azymorum quando pascha immolabant dicunt ei passover?

Biblia Sacra Vulgata

discipuli quo uis eamus et paremus tibi ut manduces pascha

14:13 And he sendeth forth two of his disciples, and saith 14:13 et mittit duos ex discipulis suis et dicit eis ite in ciuita-

unto them, Go ye into the city, and there shall meet you a man bearing a pitcher of water: follow him.

tem et occurret uobis homo laguenam aquae baiulans sequi-

14:14 And wheresoever he shall go in, say ye to the mini eum 14:14 et quocumque introierit dicite domino do-

goodman of the house, The Master saith, Where is the guestchamber, where I shall eat the passover with my dis-

ciples? mus quia magister dicit ubi est refectio mea ubi pascha cum 14:15 And he will shew you a large upper room furnished discipulis meis manducem 14:15 et ipse uobis demonstra-

and prepared: there make ready for us.

14:16 And his disciples went forth, and came into the city, bit cenaculum grande stratum et illic parate nobis 14:16 et

and found as he had said unto them: and they made ready the passover.

abierunt discipuli eius et uenerunt in ciuitatem et inuenerunt

14:17 And in the evening he cometh with the twelve. sicut dixerat illis et praeparauerunt pascha 14:17 uespere

14:18 And as they sat and did eat, Jesus said, Verily I say autem facto uenit cum duodecim 14:18 et discumbentibus unto you, One of you which eateth with me shall betray me.

eis et manducantibus ait Iesus amen dico uobis quia unus ex

14:19 And they began to be sorrowful, and to say unto him uobis me tradet qui manducat mecum 14:19 at illi coepeone by one, Is it I? and another said, Is it I?

14:20 And he answered and said unto them, It is one of runt contristari et dicere ei singillatim numquid ego 14:20

the twelve, that dippeth with me in the dish.

qui ait illis unus ex duodecim qui intinguit mecum in catino

14:21 The Son of man indeed goeth, as it is written of him: 14:21 et Filius quidem hominis uadit sicut scriptum est de

but woe to that man by whom the Son of man is betrayed! good were it for that man if he had never been born.

eo uae autem homini illi per quem Filius hominis traditur

14:22 And as they did eat, Jesus took bread, and blessed, bonum ei si non esset natus homo ille 14:22 et manducanand brake it, and gave to them, and said, Take, eat: this is my body.

tibus illis accepit Iesus panem et benedicens fregit et dedit

14:23 And he took the cup, and when he had given thanks, eis et ait sumite hoc est corpus meum 14:23 et accepto ca-

he gave it to them: and they all drank of it.

14:24 And he said unto them, This is my blood of the new lice gratias agens dedit eis et biberunt ex illo omnes 14:24

testament, which is shed for many.

et ait illis hic est sanguis meus noui testamenti qui pro mul-

14:25 Verily I say unto you, I will drink no more of the tis effunditur 14:25 amen dico uobis quod iam non bibam

fruit of the vine, until that day that I drink it new in the kingdom of God.

de genimine uitis usque in diem illum cum illud bibam no-

14:26 And when they had sung an hymn, they went out uum in regno Dei 14:26 et hymno dicto exierunt in montem

into the mount of Olives.

14:27 And Jesus saith unto them, All ye shall be offended Oliuarum 14:27 et ait eis Iesus omnes scandalizabimini in

because of me this night: for it is written, I will smite the shepherd, and the sheep shall be scattered.

nocte ista quia scriptum est percutiam pastorem et disper-

14:28 But after that I am risen, I will go before you into gentur oues 14:28 sed posteaquam resurrexero praecedam

Galilee.

14:29 But Peter said unto him, Although all shall be of- uos in Galilaeam 14:29 Petrus autem ait ei et si omnes scan-

fended, yet will not I.

14:30 And Jesus saith unto him, Verily I say unto thee, dalizati fuerint sed non ego 14:30 et ait illi Iesus amen dico

That this day, even in this night, before the cock crow twice, thou shalt deny me thrice.

tibi quia tu hodie in nocte hac priusquam bis gallus uocem

14:31 But he spake the more vehemently, If I should die dederit ter me es negaturus 14:31 at ille amplius loquebatur

with thee, I will not deny thee in any wise. Likewise also said they all.

et si oportuerit me simul conmori tibi non te negabo simili-

14:32 And they came to a place which was named Geth- ter autem et omnes dicebant 14:32 et ueniunt in praedium

semane: and he saith to his disciples, Sit ye here, while I shall pray.

Euangelium Secundum Marcum

cui nomen Gethsemani et ait discipulis suis sedete hic donec orem 14:33 et adsumit Petrum et Iacobum et Iohannem se- 14:33 And he taketh with him Peter and James and John,

and began to be sore amazed, and to be very heavy;

cum et coepit pauere et taedere 14:34 et ait illis tristis est 14:34 And saith unto them, My soul is exceeding sorrow-

ful unto death: tarry ye here, and watch.

anima mea usque ad mortem sustinete hic et uigilate 14:35 14:35 And he went forward a little, and fell on the ground,

and prayed that, if it were possible, the hour might pass from him.

et cum processisset paululum procidit super terram et orabat ut si fieri posset transiret ab eo hora 14:36 et dixit Abba 14:36 And he said, Abba, Father, all things are possible

unto thee; take away this cup from me: nevertheless not what I will, but what thou wilt.

Pater omnia possibilia tibi sunt transfer calicem hunc a me sed non quod ego uolo sed quod tu 14:37 et uenit et inuenit 14:37 And he cometh, and findeth them sleeping, and saith

unto Peter, Simon, sleepest thou? couldest not thou watch one hour?

eos dormientes et ait Petro Simon dormis non potuisti una hora uigilare 14:38 uigilate et orate ut non intretis in temp- 14:38 Watch ye and pray, lest ye enter into temptation.

The spirit truly is ready, but the flesh is weak.

tationem spiritus quidem promptus caro uero infirma 14:39 14:39 And again he went away, and prayed, and spake the

same words.

et iterum abiens orauit eundem sermonem dicens 14:40 et 14:40 And when he returned, he found them asleep again,

(for their eyes were heavy,) neither wist they what to answer him.

reuersus denuo inuenit eos dormientes erant enim oculi illorum ingrauati et ignorabant quid responderent ei 14:41 et 14:41 And he cometh the third time, and saith unto them,

Sleep on now, and take your rest: it is enough, the hour is come; behold, the Son of man is betrayed into the hands

uenit tertio et ait illis dormite iam et requiescite sufficit ue- of sinners. nit hora ecce traditur Filius hominis in manus peccatorum

14:42 surgite eamus ecce qui me tradit prope est 14:43 et 14:42 Rise up, let us go; lo, he that betrayeth me is at hand. 14:43 And immediately, while he yet spake, cometh Judas, one of the twelve, and with him a great multitude with swords and staves, from the chief priests and the scribes and the elders.

adhuc eo loquente uenit Iudas Scarioth unus ex duodecim et cum illo turba cum gladiis et lignis a summis sacerdotibus et a scribis et a senioribus 14:44 dederat autem traditor eius 14:44 And he that betrayed him had given them a token,

saying, Whomsoever I shall kiss, that same is he; take him, and lead him away safely.

signum eis dicens quemcumque osculatus fuero ipse est tenete eum et ducite 14:45 et cum uenisset statim accedens 14:45 And as soon as he was come, he goeth straightway

to him, and saith, Master, master; and kissed him.

ad eum ait rabbi et osculatus est eum 14:46 at illi manus 14:46 And they laid their hands on him, and took him. iniecerunt in eum et tenuerunt eum 14:47 unus autem qui- 14:47 And one of them that stood by drew a sword, and

smote a servant of the high priest, and cut off his ear.

dam de circumstantibus educens gladium percussit seruum summi sacerdotis et amputauit illi auriculam 14:48 et re- 14:48 And Jesus answered and said unto them, Are ye

come out, as against a thief, with swords and with staves to take me?

spondens Iesus ait illis tamquam ad latronem existis cum gladiis et lignis conprehendere me 14:49 cotidie eram apud 14:49 I was daily with you in the temple teaching, and ye

took me not: but the scriptures must be fulfilled.

uos in templo docens et non me tenuistis sed ut adimpleantur scripturae 14:50 tunc discipuli eius relinquentes eum 14:50 And they all forsook him, and fled. omnes fugerunt 14:51 adulescens autem quidam sequebatur 14:51 And there followed him a certain young man, hav-

ing a linen cloth cast about his naked body; and the young men laid hold on him:

illum amictus sindone super nudo et tenuerunt eum 14:52

14:52 And he left the linen cloth, and fled from them naked.

at ille reiecta sindone nudus profugit ab eis 14:53 et ad- 14:53 And they led Jesus away to the high priest: and with

him were assembled all the chief priests and the elders and the scribes.

Biblia Sacra Vulgata

duxerunt Iesum ad summum sacerdotem et conueniunt om-

14:54 And Peter followed him afar off, even into the nes sacerdotes et scribae et seniores 14:54 Petrus autem a

palace of the high priest: and he sat with the servants, and warmed himself at the fire.

longe secutus est eum usque intro in atrium summi sacerdotis et sedebat cum ministris et calefaciebat se ad ignem

14:55 And the chief priests and all the council sought for 14:55 summi uero sacerdotes et omne concilium quaerebant

witness against Jesus to put him to death; and found none.

aduersum Iesum testimonium ut eum morti traderent nec in-

14:56 For many bare false witness against him, but their ueniebant 14:56 multi enim testimonium falsum dicebant

witness agreed not together.

14:57 And there arose certain, and bare false witness aduersus eum et conuenientia testimonia non erant 14:57

against him, saying,

et quidam surgentes falsum testimonium ferebant aduersus

14:58 We heard him say, I will destroy this temple that is eum dicentes 14:58 quoniam nos audiuimus eum dicentem made with hands, and within three days I will build another made without hands.

ego dissoluam templum hoc manufactum et per triduum al-

14:59 But neither so did their witness agree together. iud non manufactum aedificabo 14:59 et non erat conue-

14:60 And the high priest stood up in the midst, and asked niens testimonium illorum 14:60 et exsurgens summus sa-

Jesus, saying, Answerest thou nothing? what is it which these witness against thee?

cerdos in medium interrogauit Iesum dicens non respondes

14:61 But he held his peace, and answered nothing. Again quicquam ad ea quae tibi obiciuntur ab his 14:61 ille autem

the high priest asked him, and said unto him, Art thou the Christ, the Son of the Blessed?

tacebat et nihil respondit rursum summus sacerdos interro-

14:62 And Jesus said, I am: and ye shall see the Son of gabat eum et dicit ei tu es Christus Filius Benedicti 14:62

man sitting on the right hand of power, and coming in the clouds of heaven.

Iesus autem dixit illi ego sum et uidebitis Filium hominis a dextris sedentem Virtutis et uenientem cum nubibus caeli

14:63 Then the high priest rent his clothes, and saith, What 14:63 summus autem sacerdos scindens uestimenta sua ait need we any further witnesses?

14:64 Ye have heard the blasphemy: what think ye? And quid adhuc desideramus testes 14:64 audistis blasphemiam

they all condemned him to be guilty of death.

quid uobis uidetur qui omnes condemnauerunt eum esse re-

14:65 And some began to spit on him, and to cover his um mortis 14:65 et coeperunt quidam conspuere eum et ue-

face, and to buffet him, and to say unto him, Prophesy: and the servants did strike him with the palms of their hands.

lare faciem eius et colaphis eum caedere et dicere ei pro-

14:66 And as Peter was beneath in the palace, there phetiza et ministri alapis eum caedebant 14:66 et cum esset

cometh one of the maids of the high priest:

Petrus in atrio deorsum uenit una ex ancillis summi sacerdo-

14:67 And when she saw Peter warming himself, she tis 14:67 et cum uidisset Petrum calefacientem se aspiciens

looked upon him, and said, And thou also wast with Jesus of Nazareth.

illum ait et tu cum Iesu Nazareno eras 14:68 at ille negauit

14:68 But he denied, saying, I know not, neither understand I what thou sayest. And he went out into the porch; and the cock crew.

dicens neque scio neque noui quid dicas et exiit foras ante

14:69 And a maid saw him again, and began to say to them atrium et gallus cantauit 14:69 rursus autem cum uidisset

that stood by, This is one of them.

illum ancilla coepit dicere circumstantibus quia hic ex illis

14:70 And he denied it again. And a little after, they that est 14:70 at ille iterum negauit et post pusillum rursus qui

stood by said again to Peter, Surely thou art one of them: for thou art a Galilaean, and thy speech agreeth thereto.

adstabant dicebant Petro uere ex illis es nam et Galilaeus es

Euangelium Secundum Marcum

14:71 ille autem coepit anathematizare et iurare quia nescio 14:71 But he began to curse and to swear, saying, I know

not this man of whom ye speak.

hominem istum quem dicitis 14:72 et statim iterum gallus 14:72 And the second time the cock crew. And Peter

called to mind the word that Jesus said unto him, Before the cock crow twice, thou shalt deny me thrice. And when

cantauit et recordatus est Petrus uerbi quod dixerat ei Iesus he thought thereon, he wept. priusquam gallus cantet bis ter me negabis et coepit flere

15:1 et confestim mane consilium facientes summi sacer- 15:1 And straightway in the morning the chief priests held a consultation with the elders and scribes and the whole council, and bound Jesus, and carried him away, and de-

dotes cum senioribus et scribis et uniuerso concilio uincien- livered him to Pilate. tes Iesum duxerunt et tradiderunt Pilato 15:2 et interroga- 15:2 And Pilate asked him, Art thou the King of the Jews?

And he answering said unto them, Thou sayest it.

uit eum Pilatus tu es rex Iudaeorum at ille respondens ait illi tu dicis 15:3 et accusabant eum summi sacerdotes in 15:3 And the chief priests accused him of many things:

but he answered nothing.

multis 15:4 Pilatus autem rursum interrogauit eum dicens 15:4 And Pilate asked him again, saying, Answerest thou

nothing? behold how many things they witness against thee.

non respondes quicquam uide in quantis te accusant 15:5

15:5 But Jesus yet answered nothing; so that Pilate marvelled.

Iesus autem amplius nihil respondit ita ut miraretur Pilatus 15:6 per diem autem festum dimittere solebat illis unum 15:6 Now at that feast he released unto them one prisoner,

whomsoever they desired.

ex uinctis quemcumque petissent 15:7 erat autem qui di- 15:7 And there was one named Barabbas, which lay bound

with them that had made insurrection with him, who had committed murder in the insurrection.

cebatur Barabbas qui cum seditiosis erat uinctus qui in seditione fecerant homicidium 15:8 et cum ascendisset turba 15:8 And the multitude crying aloud began to desire him

to do as he had ever done unto them.

coepit rogare sicut semper faciebat illis 15:9 Pilatus autem 15:9 But Pilate answered them, saying, Will ye that I re-

lease unto you the King of the Jews?

respondit eis et dixit uultis dimittam uobis regem Iudaeorum 15:10 sciebat enim quod per inuidiam tradidissent eum 15:10 For he knew that the chief priests had delivered him

for envy.

summi sacerdotes 15:11 pontifices autem concitauerunt tur- 15:11 But the chief priests moved the people, that he

should rather release Barabbas unto them.

bam ut magis Barabban dimitteret eis 15:12 Pilatus autem 15:12 And Pilate answered and said again unto them,

What will ye then that I shall do unto him whom ye call the King of the Jews?

iterum respondens ait illis quid ergo uultis faciam regi Iudaeorum 15:13 at illi iterum clamauerunt crucifige eum 15:14 15:13 And they cried out again, Crucify him.

15:14 Then Pilate said unto them, Why, what evil hath he done? And they cried out the more exceedingly, Crucify him.

Pilatus uero dicebat eis quid enim mali fecit at illi magis cla-

mabant crucifige eum 15:15 Pilatus autem uolens populo 15:15 And so Pilate, willing to content the people, released Barabbas unto them, and delivered Jesus, when he had scourged him, to be crucified.

satisfacere dimisit illis Barabban et tradidit Iesum flagellis caesum ut crucifigeretur 15:16 milites autem duxerunt eum 15:16 And the soldiers led him away into the hall, called

Praetorium; and they call together the whole band.

intro in atrium praetorii et conuocant totam cohortem 15:17 15:17 And they clothed him with purple, and platted a

crown of thorns, and put it about his head,

et induunt eum purpuram et inponunt ei plectentes spineam coronam 15:18 et coeperunt salutare eum haue rex Iudae- 15:18 And began to salute him, Hail, King of the Jews! orum 15:19 et percutiebant caput eius harundine et con- 15:19 And they smote him on the head with a reed, and did

spit upon him, and bowing their knees worshipped him.

spuebant eum et ponentes genua adorabant eum 15:20 et 15:20 And when they had mocked him, they took off the

purple from him, and put his own clothes on him, and led him out to crucify him.

postquam inluserunt ei exuerunt illum purpuram et indue-

Biblia Sacra Vulgata

runt eum uestimentis suis et educunt illum ut crucifigerent

15:21 And they compel one Simon a Cyrenian, who eum 15:21 et angariauerunt praetereuntem quempiam Si-

passed by, coming out of the country, the father of Alexander and Rufus, to bear his cross.

monem Cyreneum uenientem de uilla patrem Alexandri et

15:22 And they bring him unto the place Golgotha, which Rufi ut tolleret crucem eius 15:22 et perducunt illum in Gol-

is, being interpreted, The place of a skull.

15:23 And they gave him to drink wine mingled with gotha locum quod est interpretatum Caluariae locus 15:23

myrrh: but he received it not.

15:24 And when they had crucified him, they parted his et dabant ei bibere murratum uinum et non accepit 15:24 et

garments, casting lots upon them, what every man should take.

crucifigentes eum diuiserunt uestimenta eius mittentes sor-

15:25 And it was the third hour, and they crucified him. tem super eis quis quid tolleret 15:25 erat autem hora tertia 15:26 And the superscription of his accusation was written et crucifixerunt eum 15:26 et erat titulus causae eius inscri-

over, THE KING OF THE JEWS.

15:27 And with him they crucify two thieves; the one on ptus rex Iudaeorum 15:27 et cum eo crucifigunt duos lat-

his right hand, and the other on his left. 15:28 And the scripture was fulfilled, which saith, And he rones unum a dextris et alium a sinistris eius 15:28 et adwas numbered with the transgressors.

impleta est scriptura quae dicit et cum iniquis reputatus est

15:29 And they that passed by railed on him, wagging 15:29 et praetereuntes blasphemabant eum mouentes capita

their heads, and saying, Ah, thou that destroyest the temple, and buildest it in three days,

sua et dicentes ua qui destruit templum et in tribus diebus ae-

15:30 Save thyself, and come down from the cross. dificat 15:30 saluum fac temet ipsum descendens de cruce 15:31 Likewise also the chief priests mocking said among 15:31 similiter et summi sacerdotes ludentes ad alterutrum

themselves with the scribes, He saved others; himself he cannot save.

cum scribis dicebant alios saluos fecit se ipsum non potest

15:32 Let Christ the King of Israel descend now from the saluum facere 15:32 Christus rex Israhel descendat nunc de

cross, that we may see and believe. And they that were crucified with him reviled him.

cruce ut uideamus et credamus et qui cum eo crucifixi er-

15:33 And when the sixth hour was come, there was dark- ant conuiciabantur ei 15:33 et facta hora sexta tenebrae fac-

ness over the whole land until the ninth hour. 15:34 And at the ninth hour Jesus cried with a loud voice, tae sunt per totam terram usque in horam nonam 15:34 et saying, Eloi, Eloi, lama sabachthani? which is, being interpreted, My God, my God, why hast thou forsaken me?

hora nona exclamauit Iesus uoce magna dicens Heloi Heloi lama sabacthani quod est interpretatum Deus meus Deus

15:35 And some of them that stood by, when they heard meus ut quid dereliquisti me 15:35 et quidam de circum-

it, said, Behold, he calleth Elias.

15:36 And one ran and filled a spunge full of vinegar, and stantibus audientes dicebant ecce Heliam uocat 15:36 cur-

put it on a reed, and gave him to drink, saying, Let alone; let us see whether Elias will come to take him down.

rens autem unus et implens spongiam aceto circumponensque calamo potum dabat ei dicens sinite uideamus si ueniat

15:37 And Jesus cried with a loud voice, and gave up the Helias ad deponendum eum 15:37 Iesus autem emissa uoce

ghost.

15:38 And the veil of the temple was rent in twain from magna exspirauit 15:38 et uelum templi scissum est in duo

the top to the bottom. 15:39 And when the centurion, which stood over against a sursum usque deorsum 15:39 uidens autem centurio qui him, saw that he so cried out, and gave up the ghost, he said, Truly this man was the Son of God.

ex aduerso stabat quia sic clamans exspirasset ait uere homo

15:40 There were also women looking on afar off: among hic Filius Dei erat 15:40 erant autem et mulieres de longe

whom was Mary Magdalene, and Mary the mother of James the less and of Joses, and Salome;

aspicientes inter quas et Maria Magdalene et Maria Iacobi

Euangelium Secundum Marcum

minoris et Ioseph mater et Salome 15:41 et cum esset in 15:41 (Who also, when he was in Galilee, followed him,

and ministered unto him;) and many other women which came up with him unto Jerusalem.

Galilaea sequebantur eum et ministrabant ei et aliae multae quae simul cum eo ascenderant Hierosolyma 15:42 et cum 15:42 And now when the even was come, because it was

the preparation, that is, the day before the sabbath,

iam sero esset factum quia erat parasceue quod est ante sabbatum 15:43 uenit Ioseph ab Arimathia nobilis decurio qui 15:43 Joseph of Arimathaea, an honourable counsellor,

which also waited for the kingdom of God, came, and went in boldly unto Pilate, and craved the body of Jesus.

et ipse erat expectans regnum Dei et audacter introiit ad Pilatum et petiit corpus Iesu 15:44 Pilatus autem mirabatur si 15:44 And Pilate marvelled if he were already dead: and

calling unto him the centurion, he asked him whether he had been any while dead.

iam obisset et accersito centurione interrogauit eum si iam mortuus esset 15:45 et cum cognouisset a centurione dona- 15:45 And when he knew it of the centurion, he gave the

body to Joseph.

uit corpus Ioseph 15:46 Ioseph autem mercatus sindonem et 15:46 And he bought fine linen, and took him down, and

wrapped him in the linen, and laid him in a sepulchre which was hewn out of a rock, and rolled a stone unto

deponens eum inuoluit sindone et posuit eum in monumento the door of the sepulchre. quod erat excisum de petra et aduoluit lapidem ad ostium monumenti 15:47 Maria autem Magdalene et Maria Ioseph 15:47 And Mary Magdalene and Mary the mother of Joses

beheld where he was laid.

aspiciebant ubi poneretur

16:1 et cum transisset sabbatum Maria Magdalene et Ma- 16:1 And when the sabbath was past, Mary Magdalene,

and Mary the mother of James, and Salome, had bought sweet spices, that they might come and anoint him.

ria Iacobi et Salome emerunt aromata ut uenientes unguerent eum 16:2 et ualde mane una sabbatorum ueniunt ad mo- 16:2 And very early in the morning the first day of the

week, they came unto the sepulchre at the rising of the sun.

numentum orto iam sole 16:3 et dicebant ad inuicem quis

16:3 And they said among themselves, Who shall roll us away the stone from the door of the sepulchre?

reuoluet nobis lapidem ab ostio monumenti 16:4 et respi- 16:4 And when they looked, they saw that the stone was

rolled away: for it was very great.

cientes uident reuolutum lapidem erat quippe magnus ualde 16:5 et introeuntes in monumento uiderunt iuuenem seden- 16:5 And entering into the sepulchre, they saw a young

man sitting on the right side, clothed in a long white garment; and they were affrighted.

tem in dextris coopertum stola candida et obstipuerunt 16:6

16:6 And he saith unto them, Be not affrighted: Ye seek Jesus of Nazareth, which was crucified: he is risen; he is not here: behold the place where they laid him.

qui dicit illis nolite expauescere Iesum quaeritis Nazarenum crucifixum surrexit non est hic ecce locus ubi posuerunt eum 16:7 sed ite et dicite discipulis eius et Petro quia praecedit uos 16:7 But go your way, tell his disciples and Peter that he

goeth before you into Galilee: there shall ye see him, as he said unto you.

in Galilaeam ibi eum uidebitis sicut dixit uobis 16:8 at illae

16:8 And they went out quickly, and fled from the sepulchre; for they trembled and were amazed: neither said they any thing to any man; for they were afraid.

exeuntes fugerunt de monumento inuaserat enim eas tremor et pauor et nemini quicquam dixerunt timebant enim 16:9 16:9 Now when Jesus was risen early the first day of the

week, he appeared first to Mary Magdalene, out of whom he had cast seven devils.

surgens autem mane prima sabbati apparuit primo Mariae Magdalenae de qua eiecerat septem daemonia 16:10 illa ua- 16:10 And she went and told them that had been with him,

as they mourned and wept.

dens nuntiauit his qui cum eo fuerant lugentibus et flentibus 16:11 et illi audientes quia uiueret et uisus esset ab ea non 16:11 And they, when they had heard that he was alive,

and had been seen of her, believed not.

crediderunt 16:12 post haec autem duobus ex eis ambulan- 16:12 After that he appeared in another form unto two of

them, as they walked, and went into the country.

Biblia Sacra Vulgata

16:13 And they went and told it unto the residue: neither tibus ostensus est in alia effigie euntibus in uillam 16:13 et

believed they them.

16:14 Afterward he appeared unto the eleven as they sat at illi euntes nuntiauerunt ceteris nec illis crediderunt 16:14

meat, and upbraided them with their unbelief and hardness of heart, because they believed not them which had seen

him after he was risen. nouissime recumbentibus illis undecim apparuit et exprob-

rauit incredulitatem illorum et duritiam cordis quia his qui

16:15 And he said unto them, Go ye into all the world, and uiderant eum resurrexisse non crediderant 16:15 et dixit eis

preach the gospel to every creature.

euntes in mundum uniuersum praedicate euangelium omni

16:16 He that believeth and is baptized shall be saved; but creaturae 16:16 qui crediderit et baptizatus fuerit saluus erit

he that believeth not shall be damned.

16:17 And these signs shall follow them that believe; In qui uero non crediderit condemnabitur 16:17 signa autem

my name shall they cast out devils; they shall speak with new tongues;

eos qui crediderint haec sequentur in nomine meo daemonia

16:18 They shall take up serpents; and if they drink any eicient linguis loquentur nouis 16:18 serpentes tollent et si

deadly thing, it shall not hurt them; they shall lay hands on the sick, and they shall recover.

mortiferum quid biberint non eos nocebit super aegrotos ma-

16:19 So then after the Lord had spoken unto them, he was nus inponent et bene habebunt 16:19 et Dominus quidem received up into heaven, and sat on the right hand of God.

postquam locutus est eis adsumptus est in caelum et sedit a

16:20 And they went forth, and preached every where, the dextris Dei 16:20 illi autem profecti praedicauerunt ubique

Lord working with them, and confirming the word with signs following. Amen.

Domino cooperante et sermonem confirmante sequentibus signis

Euangelium Secundum Lucam

1:1 quoniam quidem multi conati sunt ordinare narratio- 1:1 Forasmuch as many have taken in hand to set forth in

order a declaration of those things which are most surely believed among us,

nem quae in nobis conpletae sunt rerum 1:2 sicut tradide-

1:2 Even as they delivered them unto us, which from the beginning were eyewitnesses, and ministers of the word;

runt nobis qui ab initio ipsi uiderunt et ministri fuerunt sermonis 1:3 uisum est et mihi adsecuto a principio omnibus 1:3 It seemed good to me also, having had perfect under-

standing of all things from the very first, to write unto thee in order, most excellent Theophilus,

diligenter ex ordine tibi scribere optime Theophile 1:4 ut

1:4 That thou mightest know the certainty of those things, wherein thou hast been instructed.

cognoscas eorum uerborum de quibus eruditus es ueritatem 1:5 fuit in diebus Herodis regis Iudaeae sacerdos quidam no- 1:5 THERE was in the days of Herod, the king of Judaea,

a certain priest named Zacharias, of the course of Abia: and his wife was of the daughters of Aaron, and her name

mine Zaccharias de uice Abia et uxor illi de filiabus Aaron et was Elisabeth. nomen eius Elisabeth 1:6 erant autem iusti ambo ante De- 1:6 And they were both righteous before God, walking in

all the commandments and ordinances of the Lord blameless.

um incedentes in omnibus mandatis et iustificationibus Domini sine querella 1:7 et non erat illis filius eo quod esset 1:7 And they had no child, because that Elisabeth was bar-

ren, and they both were now well stricken in years.

Elisabeth sterilis et ambo processissent in diebus suis 1:8 1:8 And it came to pass, that while he executed the priest’s

office before God in the order of his course,

factum est autem cum sacerdotio fungeretur in ordine uicis

suae ante Deum 1:9 secundum consuetudinem sacerdotii 1:9 According to the custom of the priest’s office, his lot was to burn incense when he went into the temple of the Lord.

sorte exiit ut incensum poneret ingressus in templum Domini 1:10 et omnis multitudo erat populi orans foris hora 1:10 And the whole multitude of the people were praying

without at the time of incense.

incensi 1:11 apparuit autem illi angelus Domini stans a de- 1:11 And there appeared unto him an angel of the Lord

standing on the right side of the altar of incense.

xtris altaris incensi 1:12 et Zaccharias turbatus est uidens 1:12 And when Zacharias saw him, he was troubled, and

fear fell upon him.

et timor inruit super eum 1:13 ait autem ad illum angelus 1:13 But the angel said unto him, Fear not, Zacharias: for thy prayer is heard; and thy wife Elisabeth shall bear thee a son, and thou shalt call his name John.

ne timeas Zaccharia quoniam exaudita est deprecatio tua et uxor tua Elisabeth pariet tibi filium et uocabis nomen eius Iohannem 1:14 et erit gaudium tibi et exultatio et multi in 1:14 And thou shalt have joy and gladness; and many shall

rejoice at his birth.

natiuitate eius gaudebunt 1:15 erit enim magnus coram Do- 1:15 For he shall be great in the sight of the Lord, and shall

drink neither wine nor strong drink; and he shall be filled with the Holy Ghost, even from his mother’s womb.

mino et uinum et sicera non bibet et Spiritu Sancto replebitur adhuc ex utero matris suae 1:16 et multos filiorum Israhel 1:16 And many of the children of Israel shall he turn to the

Lord their God.

conuertet ad Dominum Deum ipsorum 1:17 et ipse praece- 1:17 And he shall go before him in the spirit and power of

Elias, to turn the hearts of the fathers to the children, and the disobedient to the wisdom of the just; to make ready a

det ante illum in spiritu et uirtute Heliae ut conuertat corda people prepared for the Lord. patrum in filios et incredibiles ad prudentiam iustorum parare Domino plebem perfectam 1:18 et dixit Zaccharias ad 1:18 And Zacharias said unto the angel, Whereby shall I

know this? for I am an old man, and my wife well stricken in years.

angelum unde hoc sciam ego enim sum senex et uxor mea processit in diebus suis 1:19 et respondens angelus dixit ei 1:19 And the angel answering said unto him, I am Gabriel,

that stand in the presence of God; and am sent to speak unto thee, and to shew thee these glad tidings.

ego sum Gabrihel qui adsto ante Deum et missus sum loqui ad te et haec tibi euangelizare 1:20 et ecce eris tacens et 1:20 And, behold, thou shalt be dumb, and not able to

speak, until the day that these things shall be performed, because thou believest not my words, which shall be fulfilled in their season.

Biblia Sacra Vulgata

non poteris loqui usque in diem quo haec fiant pro eo quod non credidisti uerbis meis quae implebuntur in tempore suo

1:21 And the people waited for Zacharias, and marvelled 1:21 et erat plebs expectans Zacchariam et mirabantur quod

that he tarried so long in the temple.

1:22 And when he came out, he could not speak unto tardaret ipse in templo 1:22 egressus autem non poterat lo-

them: and they perceived that he had seen a vision in the temple: for he beckoned unto them, and remained speech-

less. qui ad illos et cognouerunt quod uisionem uidisset in tem-1:23 And it came to pass, that, as soon as the days of his plo et ipse erat innuens illis et permansit mutus 1:23 et fac-

ministration were accomplished, he departed to his own house.

tum est ut impleti sunt dies officii eius abiit in domum suam

1:24 And after those days his wife Elisabeth conceived, 1:24 post hos autem dies concepit Elisabeth uxor eius et oc-

and hid herself five months, saying,

1:25 Thus hath the Lord dealt with me in the days wherein cultabat se mensibus quinque dicens 1:25 quia sic mihi fe-

he looked on me, to take away my reproach among men.

cit Dominus in diebus quibus respexit auferre obprobrium

1:26 And in the sixth month the angel Gabriel was sent meum inter homines 1:26 in mense autem sexto missus est

from God unto a city of Galilee, named Nazareth,

angelus Gabrihel a Deo in ciuitatem Galilaeae cui nomen

1:27 To a virgin espoused to a man whose name was Nazareth 1:27 ad uirginem desponsatam uiro cui nomen

Joseph, of the house of David; and the virgin’s name was Mary.

erat Ioseph de domo Dauid et nomen uirginis Maria 1:28

1:28 And the angel came in unto her, and said, Hail, thou that art highly favoured, the Lord is with thee: blessed art thou among women.

et ingressus angelus ad eam dixit haue gratia plena Dominus

1:29 And when she saw him, she was troubled at his say- tecum benedicta tu in mulieribus 1:29 quae cum uidisset

ing, and cast in her mind what manner of salutation this should be.

turbata est in sermone eius et cogitabat qualis esset ista sa-

1:30 And the angel said unto her, Fear not, Mary: for thou lutatio 1:30 et ait angelus ei ne timeas Maria inuenisti enim

hast found favour with God.

1:31 And, behold, thou shalt conceive in thy womb, and gratiam apud Deum 1:31 ecce concipies in utero et paries

bring forth a son, and shalt call his name JESUS. 1:32 He shall be great, and shall be called the Son of the filium et uocabis nomen eius Iesum 1:32 hic erit magnus Highest: and the Lord God shall give unto him the throne of his father David:

et Filius Altissimi uocabitur et dabit illi Dominus Deus se-

1:33 And he shall reign over the house of Jacob for ever; dem Dauid patris eius 1:33 et regnabit in domo Iacob in and of his kingdom there shall be no end.

1:34 Then said Mary unto the angel, How shall this be, aeternum et regni eius non erit finis 1:34 dixit autem Ma-

seeing I know not a man?

ria ad angelum quomodo fiet istud quoniam uirum non cog-

1:35 And the angel answered and said unto her, The Holy nosco 1:35 et respondens angelus dixit ei Spiritus Sanctus

Ghost shall come upon thee, and the power of the Highest shall overshadow thee: therefore also that holy thing

which shall be born of thee shall be called the Son of God. superueniet in te et uirtus Altissimi obumbrabit tibi ideoque

1:36 And, behold, thy cousin Elisabeth, she hath also con- et quod nascetur sanctum uocabitur Filius Dei 1:36 et ecce ceived a son in her old age: and this is the sixth month with her, who was called barren.

Elisabeth cognata tua et ipsa concepit filium in senecta sua

1:37 For with God nothing shall be impossible. et hic mensis est sextus illi quae uocatur sterilis 1:37 quia

1:38 And Mary said, Behold the handmaid of the Lord; be non erit inpossibile apud Deum omne uerbum 1:38 dixit it unto me according to thy word. And the angel departed from her.

autem Maria ecce ancilla Domini fiat mihi secundum uer-

1:39 And Mary arose in those days, and went into the hill bum tuum et discessit ab illa angelus 1:39 exsurgens autem

country with haste, into a city of Juda;

Euangelium Secundum Lucam

Maria in diebus illis abiit in montana cum festinatione in ciuitatem Iuda 1:40 et intrauit in domum Zacchariae et sa- 1:40 And entered into the house of Zacharias, and saluted

Elisabeth.

lutauit Elisabeth 1:41 et factum est ut audiuit salutationem 1:41 And it came to pass, that, when Elisabeth heard the

salutation of Mary, the babe leaped in her womb; and Elisabeth was filled with the Holy Ghost:

Mariae Elisabeth exultauit infans in utero eius et repleta est Spiritu Sancto Elisabeth 1:42 et exclamauit uoce magna et 1:42 And she spake out with a loud voice, and said,

Blessed art thou among women, and blessed is the fruit of thy womb.

dixit benedicta tu inter mulieres et benedictus fructus uentris tui 1:43 et unde hoc mihi ut ueniat mater Domini mei 1:43 And whence is this to me, that the mother of my Lord

should come to me?

ad me 1:44 ecce enim ut facta est uox salutationis tuae in 1:44 For, lo, as soon as the voice of thy salutation sounded

in mine ears, the babe leaped in my womb for joy.

auribus meis exultauit in gaudio infans in utero meo 1:45 et 1:45 And blessed is she that believed: for there shall be a

performance of those things which were told her from the Lord.

beata quae credidit quoniam perficientur ea quae dicta sunt ei a Domino 1:46 et ait Maria magnificat anima mea Do- 1:46 And Mary said, My soul doth magnify the Lord, minum 1:47 et exultauit spiritus meus in Deo salutari meo 1:47 And my spirit hath rejoiced in God my Saviour. 1:48 quia respexit humilitatem ancillae suae ecce enim ex hoc 1:48 For he hath regarded the low estate of his hand-

maiden: for, behold, from henceforth all generations shall call me blessed.

beatam me dicent omnes generationes 1:49 quia fecit mihi

1:49 For he that is mighty hath done to me great things; and holy is his name.

magna qui potens est et sanctum nomen eius 1:50 et mise- 1:50 And his mercy is on them that fear him from genera-

tion to generation.

ricordia eius in progenies et progenies timentibus eum 1:51 1:51 He hath shewed strength with his arm; he hath scat-

tered the proud in the imagination of their hearts.

fecit potentiam in brachio suo dispersit superbos mente cordis sui 1:52 deposuit potentes de sede et exaltauit humiles 1:52 He hath put down the mighty from their seats, and

exalted them of low degree.

1:53 esurientes impleuit bonis et diuites dimisit inanes 1:54 1:53 He hath filled the hungry with good things; and the

rich he hath sent empty away.

1:54 He hath holpen his servant Israel, in remembrance of his mercy;

suscepit Israhel puerum suum memorari misericordiae 1:55

1:55 As he spake to our fathers, to Abraham, and to his

sicut locutus est ad patres nostros Abraham et semini eius in _{seed for ever.} saecula 1:56 mansit autem Maria cum illa quasi mensibus 1:56 And Mary abode with her about three months, and

returned to her own house.

tribus et reuersa est in domum suam 1:57 Elisabeth autem 1:57 Now Elisabeth’s full time came that she should be

delivered; and she brought forth a son.

impletum est tempus pariendi et peperit filium 1:58 et au- 1:58 And her neighbours and her cousins heard how the

Lord had shewed great mercy upon her; and they rejoiced with her.

dierunt uicini et cognati eius quia magnificauit Dominus misericordiam suam cum illa et congratulabantur ei 1:59 et 1:59 And it came to pass, that on the eighth day they came

to circumcise the child; and they called him Zacharias, after the name of his father.

factum est in die octauo uenerunt circumcidere puerum et uocabant eum nomine patris eius Zacchariam 1:60 et res- 1:60 And his mother answered and said, Not so; but he

shall be called John.

pondens mater eius dixit nequaquam sed uocabitur Iohannes 1:61 et dixerunt ad illam quia nemo est in cognatione tua qui 1:61 And they said unto her, There is none of thy kindred

that is called by this name.

uocetur hoc nomine 1:62 innuebant autem patri eius quem 1:62 And they made signs to his father, how he would have

him called.

uellet uocari eum 1:63 et postulans pugillarem scripsit di- 1:63 And he asked for a writing table, and wrote, saying,

His name is John. And they marvelled all.

cens Iohannes est nomen eius et mirati sunt uniuersi 1:64 1:64 And his mouth was opened immediately, and his

tongue loosed, and he spake, and praised God.

apertum est autem ilico os eius et lingua eius et loquebatur

Biblia Sacra Vulgata

1:65 And fear came on all that dwelt round about them: benedicens Deum 1:65 et factus est timor super omnes ui-

and all these sayings were noised abroad throughout all the hill country of Judaea.

cinos eorum et super omnia montana Iudaeae diuulgabantur

1:66 And all they that heard them laid them up in their omnia uerba haec 1:66 et posuerunt omnes qui audierant in

hearts, saying, What manner of child shall this be! And the hand of the Lord was with him.

corde suo dicentes quid putas puer iste erit etenim manus

1:67 And his father Zacharias was filled with the Holy Domini erat cum illo 1:67 et Zaccharias pater eius imple-

Ghost, and prophesied, saying,

1:68 Blessed be the Lord God of Israel; for he hath visited tus est Spiritu Sancto et prophetauit dicens 1:68 benedictus

and redeemed his people,

Deus Israhel quia uisitauit et fecit redeeptionem plebi suae

1:69 And hath raised up an horn of salvation for us in the 1:69 et erexit cornu salutis nobis in domo Dauid pueri sui

house of his servant David;

1:70 As he spake by the mouth of his holy prophets, which 1:70 sicut locutus est per os sanctorum qui a saeculo sunt pro-

have been since the world began:

1:71 That we should be saved from our enemies, and from phetarum eius 1:71 salutem ex inimicis nostris et de manu

the hand of all that hate us;

omnium qui oderunt nos

1:72 To perform the mercy promised to our fathers, and to 1:72 ad faciendam misericordiam cum patribus nostris et me-

remember his holy covenant;

1:73 The oath which he sware to our father Abraham, morari testamenti sui sancti 1:73 iusiurandum quod iurauit

1:74 That he would grant unto us, that we being delivered ad Abraham patrem nostrum 1:74 daturum se nobis ut sine out of the hand of our enemies might serve him without fear,

timore de manu inimicorum nostrorum liberati seruiamus illi

1:75 In holiness and righteousness before him, all the days 1:75 in sanctitate et iustitia coram ipso omnibus diebus no-

of our life.

1:76 And thou, child, shalt be called the prophet of the stris 1:76 et tu puer propheta Altissimi uocaberis praeibis

Highest: for thou shalt go before the face of the Lord to prepare his ways;

enim ante faciem Domini parare uias eius 1:77 ad dandam

1:77 To give knowledge of salvation unto his people by the remission of their sins,

scientiam salutis plebi eius in remissionem peccatorum eo-

1:78 Through the tender mercy of our God; whereby the rum 1:78 per uiscera misericordiae Dei nostri in quibus ui-

dayspring from on high hath visited us,

1:79 To give light to them that sit in darkness and in the sitauit nos oriens ex alto 1:79 inluminare his qui in tene-

shadow of death, to guide our feet into the way of peace.

bris et in umbra mortis sedent ad dirigendos pedes nostros

1:80 And the child grew, and waxed strong in spirit, and in uiam pacis 1:80 puer autem crescebat et confortabatur

was in the deserts till the day of his shewing unto Israel.

spiritu et erat in deserto usque in diem ostensionis suae ad Israhel

2:1 And it came to pass in those days, that there went out 2:1 factum est autem in diebus illis exiit edictum a Cae-

a decree from Caesar Augustus that all the world should be taxed.

sare Augusto ut describeretur uniuersus orbis 2:2 haec de-

2:2 (And this taxing was first made when Cyrenius was governor of Syria.)

2:3 And all went to be taxed, every one into his own city. scriptio prima facta est praeside Syriae Cyrino 2:3 et ibant 2:4 And Joseph also went up from Galilee, out of the city omnes ut profiterentur singuli in suam ciuitatem 2:4 ascen-

of Nazareth, into Judaea, unto the city of David, which is called Bethlehem; (because he was of the house and

lineage of David:) dit autem et Ioseph a Galilaea de ciuitate Nazareth in Iu-

daeam ciuitatem Dauid quae uocatur Bethleem eo quod es-

2:5 To be taxed with Mary his espoused wife, being great set de domo et familia Dauid 2:5 ut profiteretur cum Mawith child.

2:6 And so it was, that, while they were there, the days ria desponsata sibi uxore praegnate 2:6 factum est autem

were accomplished that she should be delivered.

Euangelium Secundum Lucam

cum essent ibi impleti sunt dies ut pareret 2:7 et peperit 2:7 And she brought forth her firstborn son, and wrapped

him in swaddling clothes, and laid him in a manger; because there was no room for them in the inn.

filium suum primogenitum et pannis eum inuoluit et reclinauit eum in praesepio quia non erat eis locus in diuersorio 2:8 et pastores erant in regione eadem uigilantes et cu- 2:8 And there were in the same country shepherds abiding

in the field, keeping watch over their flock by night.

stodientes uigilias noctis supra gregem suum 2:9 et ecce 2:9 And, lo, the angel of the Lord came upon them, and

the glory of the Lord shone round about them: and they were sore afraid.

angelus Domini stetit iuxta illos et claritas Dei circumfulsit illos et timuerunt timore magno 2:10 et dixit illis angelus 2:10 And the angel said unto them, Fear not: for, behold,

I bring you good tidings of great joy, which shall be to all people.

nolite timere ecce enim euangelizo uobis gaudium magnum quod erit omni populo 2:11 quia natus est uobis hodie sal- 2:11 For unto you is born this day in the city of David a

Saviour, which is Christ the Lord.

uator qui est Christus Dominus in ciuitate Dauid 2:12 et hoc 2:12 And this shall be a sign unto you; Ye shall find the

babe wrapped in swaddling clothes, lying in a manger.

uobis signum inuenietis infantem pannis inuolutum et positum in praesepio 2:13 et subito facta est cum angelo multi- 2:13 And suddenly there was with the angel a multitude of

the heavenly host praising God, and saying,

tudo militiae caelestis laudantium Deum et dicentium 2:14 2:14 Glory to God in the highest, and on earth peace, good

will toward men.

gloria in altissimis Deo et in terra pax in hominibus bonae uoluntatis 2:15 et factum est ut discesserunt ab eis angeli 2:15 And it came to pass, as the angels were gone away

from them into heaven, the shepherds said one to another, Let us now go even unto Bethlehem, and see this thing which is come to pass, which the Lord hath made known unto us.

in caelum pastores loquebantur ad inuicem transeamus usque Bethleem et uideamus hoc uerbum quod factum est quod fecit Dominus et ostendit nobis 2:16 et uenerunt festinan- 2:16 And they came with haste, and found Mary, and

Joseph, and the babe lying in a manger.

tes et inuenerunt Mariam et Ioseph et infantem positum in praesepio 2:17 uidentes autem cognouerunt de uerbo quod 2:17 And when they had seen it, they made known abroad

the saying which was told them concerning this child.

dictum erat illis de puero hoc 2:18 et omnes qui audierunt 2:18 And all they that heard it wondered at those things

which were told them by the shepherds.

mirati sunt et de his quae dicta erant a pastoribus ad ipsos 2:19 Maria autem conseruabat omnia uerba haec conferens 2:19 But Mary kept all these things, and pondered them in

her heart.

in corde suo 2:20 et reuersi sunt pastores glorificantes et 2:20 And the shepherds returned, glorifying and praising

God for all the things that they had heard and seen, as it was told unto them.

laudantes Deum in omnibus quae audierant et uiderant sicut dictum est ad illos 2:21 et postquam consummati sunt 2:21 And when eight days were accomplished for the cir-

cumcising of the child, his name was called JESUS, which was so named of the angel before he was conceived in the

dies octo ut circumcideretur uocatum est nomen eius Iesus womb. quod uocatum est ab angelo priusquam in utero conciperetur 2:22 et postquam impleti sunt dies purgationis eius secun- 2:22 And when the days of her purification according to

the law of Moses were accomplished, they brought him to Jerusalem, to present him to the Lord;

dum legem Mosi tulerunt illum in Hierusalem ut sisterent eum Domino 2:23 sicut scriptum est in lege Domini quia 2:23 (As it is written in the law of the LORD, Every male

that openeth the womb shall be called holy to the Lord;)

omne masculinum adaperiens uuluam sanctum Domino uo-

cabitur 2:24 et ut darent hostiam secundum quod dictum 2:24 And to offer a sacrifice according to that which is said in the law of the Lord, A pair of turtledoves, or two young pigeons.

Biblia Sacra Vulgata

est in lege Domini par turturum aut duos pullos columba-

2:25 And, behold, there was a man in Jerusalem, whose rum 2:25 et ecce homo erat in Hierusalem cui nomen Syname was Simeon; and the same man was just and devout, waiting for the consolation of Israel: and the Holy Ghost

was upon him. meon et homo iste iustus et timoratus expectans consola-2:26 And it was revealed unto him by the Holy Ghost, that tionem Israhel et Spiritus Sanctus erat in eo 2:26 et res-

he should not see death, before he had seen the Lord’s Christ.

ponsum acceperat ab Spiritu Sancto non uisurum se mortem

2:27 And he came by the Spirit into the temple: and when nisi prius uideret Christum Domini 2:27 et uenit in Spiritu

the parents brought in the child Jesus, to do for him after the custom of the law,

in templum et cum inducerent puerum Iesum parentes eius

2:28 Then took he him up in his arms, and blessed God, ut facerent secundum consuetudinem legis pro eo 2:28 et

and said,

ipse accepit eum in ulnas suas et benedixit Deum et dixit

2:29 Lord, now lettest thou thy servant depart in peace, 2:29 nunc dimittis seruum tuum Domine secundum uerbum according to thy word:

2:30 For mine eyes have seen thy salvation, tuum in pace 2:30 quia uiderunt oculi mei salutare tuum

2:31 Which thou hast prepared before the face of all peo- 2:31 quod parasti ante faciem omnium populorum 2:32 luple; 2:32 A light to lighten the Gentiles, and the glory of thy people Israel.

men ad reuelationem gentium et gloriam plebis tuae Israhel

2:33 And Joseph and his mother marvelled at those things 2:33 et erat pater eius et mater mirantes super his quae di-

which were spoken of him.

2:34 And Simeon blessed them, and said unto Mary his cebantur de illo 2:34 et benedixit illis Symeon et dixit ad

mother, Behold, this child is set for the fall and rising again of many in Israel; and for a sign which shall be spoken

against; Mariam matrem eius ecce positus est hic in ruinam et resur-

rectionem multorum in Israhel et in signum cui contradicetur

2:35 (Yea, a sword shall pierce through thy own soul also,) 2:35 et tuam ipsius animam pertransiet gladius ut reuelentur

that the thoughts of many hearts may be revealed.

2:36 And there was one Anna, a prophetess, the daughter ex multis cordibus cogitationes 2:36 et erat Anna prophe-

of Phanuel, of the tribe of Aser: she was of a great age, and had lived with an husband seven years from her virginity;

tissa filia Phanuhel de tribu Aser haec processerat in diebus multis et uixerat cum uiro suo annis septem a uirginitate

2:37 And she was a widow of about fourscore and four sua 2:37 et haec uidua usque ad annos octoginta quattuor

years, which departed not from the temple, but served God with fastings and prayers night and day.

quae non discedebat de templo ieiuniis et obsecrationibus

2:38 And she coming in that instant gave thanks likewise seruiens nocte ac die 2:38 et haec ipsa hora superueniens

unto the Lord, and spake of him to all them that looked for redemption in Jerusalem.

confitebatur Domino et loquebatur de illo omnibus qui ex-

2:39 And when they had performed all things according pectabant redeeptionem Hierusalem 2:39 et ut perfecerunt

to the law of the Lord, they returned into Galilee, to their own city Nazareth.

omnia secundum legem Domini reuersi sunt in Galilaeam in

2:40 And the child grew, and waxed strong in spirit, filled ciuitatem suam Nazareth 2:40 puer autem crescebat et con-

with wisdom: and the grace of God was upon him.

2:41 Now his parents went to Jerusalem every year at the fortabatur plenus sapientia et gratia Dei erat in illo 2:41 et

feast of the passover.

ibant parentes eius per omnes annos in Hierusalem in die

2:42 And when he was twelve years old, they went up to sollemni paschae 2:42 et cum factus esset annorum duode-

Jerusalem after the custom of the feast.

cim ascendentibus illis in Hierosolymam secundum consue-

2:43 And when they had fulfilled the days, as they re- tudinem diei festi 2:43 consummatisque diebus cum red-

turned, the child Jesus tarried behind in Jerusalem; and Joseph and his mother knew not of it.

Euangelium Secundum Lucam

iient remansit puer Iesus in Hierusalem et non cognouerunt parentes eius 2:44 existimantes autem illum esse in comi- 2:44 But they, supposing him to have been in the company,

went a day’s journey; and they sought him among their kinsfolk and acquaintance.

tatu uenerunt iter diei et requirebant eum inter cognatos et notos 2:45 et non inuenientes regressi sunt in Hierusalem 2:45 And when they found him not, they turned back again

to Jerusalem, seeking him.

requirentes eum 2:46 et factum est post triduum inuenerunt 2:46 And it came to pass, that after three days they found

him in the temple, sitting in the midst of the doctors, both hearing them, and asking them questions.

illum in templo sedentem in medio doctorum audientem illos et interrogantem 2:47 stupebant autem omnes qui eum 2:47 And all that heard him were astonished at his under-

standing and answers.

audiebant super prudentia et responsis eius 2:48 et uidentes 2:48 And when they saw him, they were amazed: and his

mother said unto him, Son, why hast thou thus dealt with us? behold, thy father and I have sought thee sorrowing.

admirati sunt et dixit mater eius ad illum fili quid fecisti nobis sic ecce pater tuus et ego dolentes quaerebamus te 2:49 2:49 And he said unto them, How is it that ye sought me?

wist ye not that I must be about my Father’s business?

et ait ad illos quid est quod me quaerebatis nesciebatis quia in his quae Patris mei sunt oportet me esse 2:50 et ipsi non 2:50 And they understood not the saying which he spake

unto them.

intellexerunt uerbum quod locutus est ad illos 2:51 et de- 2:51 And he went down with them, and came to Nazareth,

and was subject unto them: but his mother kept all these sayings in her heart.

scendit cum eis et uenit Nazareth et erat subditus illis et mater eius conseruabat omnia uerba haec in corde suo 2:52 et 2:52 And Jesus increased in wisdom and stature, and in

favour with God and man.

Iesus proficiebat sapientia aetate et gratia apud Deum et homines

3:1 anno autem quintodecimo imperii Tiberii Caesaris pro- 3:1 Now in the fifteenth year of the reign of Tiberius Cae-

sar, Pontius Pilate being governor of Judaea, and Herod being tetrarch of Galilee, and his brother Philip tetrarch of Ituraea and of the region of Trachonitis, and Lysanias the tetrarch of Abilene,

curante Pontio Pilato Iudaeam tetrarcha autem Galilaeae Herode Philippo autem fratre eius tetrarcha Itureae et Trachonitidis regionis et Lysania Abilinae tetrarcha 3:2 sub prin- 3:2 Annas and Caiaphas being the high priests, the word

of God came unto John the son of Zacharias in the wilderness.

cipibus sacerdotum Anna et Caiapha factum est uerbum Dei

super Iohannem Zacchariae filium in deserto 3:3 et uenit 3:3 And he came into all the country about Jordan, preach-

ing the baptism of repentance for the remission of sins;

in omnem regionem Iordanis praedicans baptismum paenitentiae in remissionem peccatorum 3:4 sicut scriptum est 3:4 As it is written in the book of the words of Esaias the

prophet, saying, The voice of one crying in the wilderness, Prepare ye the way of the Lord, make his paths straight.

in libro sermonum Esaiae prophetae uox clamantis in deserto parate uiam Domini rectas facite semitas eius 3:5 om- 3:5 Every valley shall be filled, and every mountain and

hill shall be brought low; and the crooked shall be made straight, and the rough ways shall be made smooth;

nis uallis implebitur et omnis mons et collis humiliabitur et erunt praua in directa et aspera in uias planas 3:6 et uide- 3:6 And all flesh shall see the salvation of God. bit omnis caro salutare Dei 3:7 dicebat ergo ad turbas quae 3:7 Then said he to the multitude that came forth to be

baptized of him, O generation of vipers, who hath warned you to flee from the wrath to come?

exiebant ut baptizarentur ab ipso genimina uiperarum quis ostendit uobis fugere a uentura ira 3:8 facite ergo fructus 3:8 Bring forth therefore fruits worthy of repentance, and

begin not to say within yourselves, We have Abraham to our father: for I say unto you, That God is able of these

dignos paenitentiae et ne coeperitis dicere patrem habemus stones to raise up children unto Abraham.

Biblia Sacra Vulgata

Abraham dico enim uobis quia potest Deus de lapidibus is-

3:9 And now also the axe is laid unto the root of the trees: tis suscitare filios Abrahae 3:9 iam enim securis ad radi-

every tree therefore which bringeth not forth good fruit is hewn down, and cast into the fire.

cem arborum posita est omnis ergo arbor non faciens fruc-

3:10 And the people asked him, saying, What shall we do tum exciditur et in ignem mittitur 3:10 et interrogabant eum

then?

3:11 He answereth and saith unto them, He that hath two turbae dicentes quid ergo faciemus 3:11 respondens autem

coats, let him impart to him that hath none; and he that hath meat, let him do likewise.

dicebat illis qui habet duas tunicas det non habenti et qui ha-

3:12 Then came also publicans to be baptized, and said bet escas similiter faciat 3:12 uenerunt autem et publicani

unto him, Master, what shall we do?

ut baptizarentur et dixerunt ad illum magister quid faciemus

3:13 And he said unto them, Exact no more than that 3:13 at ille dixit ad eos nihil amplius quam quod constitutum

which is appointed you.

3:14 And the soldiers likewise demanded of him, saying, est uobis faciatis 3:14 interrogabant autem eum et milites

And what shall we do? And he said unto them, Do violence to no man, neither accuse any falsely; and be content

with your wages. dicentes quid faciemus et nos et ait illis neminem concu-

tiatis neque calumniam faciatis et contenti estote stipendiis

3:15 And as the people were in expectation, and all men uestris 3:15 existimante autem populo et cogitantibus om-

mused in their hearts of John, whether he were the Christ, or not;

nibus in cordibus suis de Iohanne ne forte ipse esset Chri-

3:16 John answered, saying unto them all, I indeed baptize stus 3:16 respondit Iohannes dicens omnibus ego quidem you with water; but one mightier than I cometh, the latchet of whose shoes I am not worthy to unloose: he shall bap-

tize you with the Holy Ghost and with fire: aqua baptizo uos uenit autem fortior me cuius non sum di-

gnus soluere corrigiam calciamentorum eius ipse uos bap-

3:17 Whose fan is in his hand, and he will throughly purge tizabit in Spiritu Sancto et igni 3:17 cuius uentilabrum in

his floor, and will gather the wheat into his garner; but the chaff he will burn with fire unquenchable.

manu eius et purgabit aream suam et congregabit triticum in horreum suum paleas autem conburet igni inextinguibili

3:18 And many other things in his exhortation preached he 3:18 multa quidem et alia exhortans euangelizabat populum

unto the people.

3:19 But Herod the tetrarch, being reproved by him for 3:19 Herodes autem tetrarcha cum corriperetur ab illo de He-

Herodias his brother Philip’s wife, and for all the evils which Herod had done,

rodiade uxore fratris sui et de omnibus malis quae fecit He-

3:20 Added yet this above all, that he shut up John in rodes 3:20 adiecit et hoc supra omnia et inclusit Iohannem

prison.

3:21 Now when all the people were baptized, it came in carcere 3:21 factum est autem cum baptizaretur omnis

to pass, that Jesus also being baptized, and praying, the heaven was opened,

populus et Iesu baptizato et orante apertum est caelum 3:22

3:22 And the Holy Ghost descended in a bodily shape like a dove upon him, and a voice came from heaven, which said, Thou art my beloved Son; in thee I am well pleased.

et descendit Spiritus Sanctus corporali specie sicut columba in ipsum et uox de caelo facta est tu es Filius meus dilectus in

3:23 And Jesus himself began to be about thirty years of te conplacuit mihi 3:23 et ipse Iesus erat incipiens quasi an-

age, being (as was supposed) the son of Joseph, which was the son of Heli,

norum triginta ut putabatur filius Ioseph qui fuit Heli 3:24

3:24 Which was the son of Matthat, which was the son of Levi, which was the son of Melchi, which was the son of Janna, which was the son of Joseph,

qui fuit Matthat qui fuit Leui qui fuit Melchi qui fuit Iannae

3:25 Which was the son of Mattathias, which was the son qui fuit Ioseph 3:25 qui fuit Matthathiae qui fuit Amos qui

of Amos, which was the son of Naum, which was the son of Esli, which was the son of Nagge,

fuit Naum qui fuit Esli qui fuit Naggae 3:26 qui fuit Maath

3:26 Which was the son of Maath, which was the son of Mattathias, which was the son of Semei, which was the son of Joseph, which was the son of Juda,

Euangelium Secundum Lucam

qui fuit Matthathiae qui fuit Semei qui fuit Iosech qui fuit Ioda 3:27 qui fuit Iohanna qui fuit Resa qui fuit Zorobabel 3:27 Which was the son of Joanna, which was the son of

Rhesa, which was the son of Zorobabel, which was the son of Salathiel, which was the son of Neri,

qui fuit Salathihel qui fuit Neri 3:28 qui fuit Melchi qui fuit

3:28 Which was the son of Melchi, which was the son of Addi, which was the son of Cosam, which was the son of Elmodam, which was the son of Er,

Addi qui fuit Cosam qui fuit Helmadam qui fuit Her 3:29

3:29 Which was the son of Jose, which was the son of Eliezer, which was the son of Jorim, which was the son of

qui fuit Iesu qui fuit Eliezer qui fuit Iorim qui fuit Matthat _{Matthat, which was the son of Levi,} qui fuit Leui 3:30 qui fuit Symeon qui fuit Iuda qui fuit Io- 3:30 Which was the son of Simeon, which was the son of

Juda, which was the son of Joseph, which was the son of Jonan, which was the son of Eliakim,

seph qui fuit Iona qui fuit Eliachim 3:31 qui fuit Melea qui

3:31 Which was the son of Melea, which was the son of Menan, which was the son of Mattatha, which was the son of Nathan, which was the son of David,

fuit Menna qui fuit Matthata qui fuit Nathan qui fuit Dauid 3:32 qui fuit Iesse qui fuit Obed qui fuit Booz qui fuit Salmon 3:32 Which was the son of Jesse, which was the son of

Obed, which was the son of Booz, which was the son of Salmon, which was the son of Naasson,

qui fuit Naasson 3:33 qui fuit Aminadab qui fuit Aram qui

3:33 Which was the son of Aminadab, which was the son of Aram, which was the son of Esrom, which was the son of Phares, which was the son of Juda,

fuit Esrom qui fuit Phares qui fuit Iudae 3:34 qui fuit Iacob

3:34 Which was the son of Jacob, which was the son of Isaac, which was the son of Abraham, which was the son

qui fuit Isaac qui fuit Abraham qui fuit Thare qui fuit Na- _{of Thara, which was the son of Nachor,} chor 3:35 qui fuit Seruch qui fuit Ragau qui fuit Phalec qui 3:35 Which was the son of Saruch, which was the son of

Ragau, which was the son of Phalec, which was the son of Heber, which was the son of Sala,

fuit Eber qui fuit Sale 3:36 qui fuit Cainan qui fuit Arfaxat

3:36 Which was the son of Cainan, which was the son of Arphaxad, which was the son of Sem, which was the son of Noe, which was the son of Lamech,

qui fuit Sem qui fuit Noe qui fuit Lamech 3:37 qui fuit Ma-

3:37 Which was the son of Mathusala, which was the son of Enoch, which was the son of Jared, which was the son

thusalae qui fuit Enoch qui fuit Iared qui fuit Malelehel qui _{of Maleleel, which was the son of Cainan,} fuit Cainan 3:38 qui fuit Enos qui fuit Seth qui fuit Adam 3:38 Which was the son of Enos, which was the son of

Seth, which was the son of Adam, which was the son of God.

qui fuit Dei

4:1 Iesus autem plenus Spiritu Sancto regressus est ab Ior- 4:1 And Jesus being full of the Holy Ghost returned from

Jordan, and was led by the Spirit into the wilderness,

dane et agebatur in Spiritu in desertum 4:2 diebus quadra- 4:2 Being forty days tempted of the devil. And in those

days he did eat nothing: and when they were ended, he afterward hungered.

ginta et temptabatur a diabolo et nihil manducauit in diebus illis et consummatis illis esuriit 4:3 dixit autem illi diabolus 4:3 And the devil said unto him, If thou be the Son of God,

command this stone that it be made bread.

si Filius Dei es dic lapidi huic ut panis fiat 4:4 et respondit 4:4 And Jesus answered him, saying, It is written, That

man shall not live by bread alone, but by every word of God.

ad illum Iesus scriptum est quia non in pane solo uiuet homo sed in omni uerbo Dei 4:5 et duxit illum diabolus et osten- 4:5 And the devil, taking him up into an high mountain,

shewed unto him all the kingdoms of the world in a moment of time.

dit illi omnia regna orbis terrae in momento temporis 4:6 et

4:6 And the devil said unto him, All this power will I give thee, and the glory of them: for that is delivered unto me; and to whomsoever I will I give it.

ait ei tibi dabo potestatem hanc uniuersam et gloriam illorum quia mihi tradita sunt et cui uolo do illa 4:7 tu ergo si ad- 4:7 If thou therefore wilt worship me, all shall be thine. oraueris coram me erunt tua omnia 4:8 et respondens Iesus 4:8 And Jesus answered and said unto him, Get thee be-

hind me, Satan: for it is written, Thou shalt worship the Lord thy God, and him only shalt thou serve.

dixit illi scriptum est Dominum Deum tuum adorabis et illi soli seruies 4:9 et duxit illum in Hierusalem et statuit eum 4:9 And he brought him to Jerusalem, and set him on a

pinnacle of the temple, and said unto him, If thou be the Son of God, cast thyself down from hence:

supra pinnam templi et dixit illi si Filius Dei es mitte te hinc

deorsum 4:10 scriptum est enim quod angelis suis manda- 4:10 For it is written, He shall give his angels charge over thee, to keep thee:

Biblia Sacra Vulgata

4:11 And in their hands they shall bear thee up, lest at any bit de te ut conseruent te 4:11 et quia in manibus tollent te

time thou dash thy foot against a stone. 4:12 And Jesus answering said unto him, It is said, Thou ne forte offendas ad lapidem pedem tuum 4:12 et responshalt not tempt the Lord thy God.

dens Iesus ait illi dictum est non temptabis Dominum Deum

4:13 And when the devil had ended all the temptation, he tuum 4:13 et consummata omni temptatione diabolus redeparted from him for a season.

4:14 And Jesus returned in the power of the Spirit into cessit ab illo usque ad tempus 4:14 et regressus est Iesus in

Galilee: and there went out a fame of him through all the region round about.

uirtute Spiritus in Galilaeam et fama exiit per uniuersam re-

4:15 And he taught in their synagogues, being glorified of gionem de illo 4:15 et ipse docebat in synagogis eorum et all.

4:16 And he came to Nazareth, where he had been brought magnificabatur ab omnibus 4:16 et uenit Nazareth ubi erat

up: and, as his custom was, he went into the synagogue on the sabbath day, and stood up for to read.

nutritus et intrauit secundum consuetudinem suam die sab-

4:17 And there was delivered unto him the book of the bati in synagogam et surrexit legere 4:17 et traditus est illi

prophet Esaias. And when he had opened the book, he found the place where it was written,

liber prophetae Esaiae et ut reuoluit librum inuenit locum ubi

4:18 The Spirit of the Lord is upon me, because he hath scriptum erat 4:18 Spiritus Domini super me propter quod

anointed me to preach the gospel to the poor; he hath sent me to heal the brokenhearted, to preach deliverance to the captives, and recovering of sight to the blind, to set at liberty them that are bruised,

unxit me euangelizare pauperibus misit me 4:19 praedicare

4:19 To preach the acceptable year of the Lord.

captiuis remissionem et caecis uisum dimittere confractos in remissionem praedicare annum Domini acceptum et diem

4:20 And he closed the book, and he gave it again to the retributionis 4:20 et cum plicuisset librum reddidit ministro

minister, and sat down. And the eyes of all them that were in the synagogue were fastened on him.

et sedit et omnium in synagoga oculi erant intendentes in

4:21 And he began to say unto them, This day is this scrip- eum 4:21 coepit autem dicere ad illos quia hodie impleta ture fulfilled in your ears.

4:22 And all bare him witness, and wondered at the gra- est haec scriptura in auribus uestris 4:22 et omnes testimo-

cious words which proceeded out of his mouth. And they said, Is not this Joseph’s son?

nium illi dabant et mirabantur in uerbis gratiae quae procedebant de ore ipsius et dicebant nonne hic filius est Ioseph

4:23 And he said unto them, Ye will surely say unto me 4:23 et ait illis utique dicetis mihi hanc similitudinem medice

this proverb, Physician, heal thyself: whatsoever we have heard done in Capernaum, do also here in thy country.

cura te ipsum quanta audiuimus facta in Capharnaum fac et

4:24 And he said, Verily I say unto you, No prophet is hic in patria tua 4:24 ait autem amen dico uobis quia nemo

accepted in his own country.

4:25 But I tell you of a truth, many widows were in Israel propheta acceptus est in patria sua 4:25 in ueritate dico uo-

in the days of Elias, when the heaven was shut up three years and six months, when great famine was throughout

all the land; bis multae uiduae erant in diebus Heliae in Israhel quando

clusum est caelum annis tribus et mensibus sex cum facta

4:26 But unto none of them was Elias sent, save unto est fames magna in omni terra 4:26 et ad nullam illarum

Sarepta, a city of Sidon, unto a woman that was a widow.

missus est Helias nisi in Sareptha Sidoniae ad mulierem ui-

4:27 And many lepers were in Israel in the time of Eliseus duam 4:27 et multi leprosi erant in Israhel sub Heliseo pro-

the prophet; and none of them was cleansed, saving Naaman the Syrian.

pheta et nemo eorum mundatus est nisi Neman Syrus 4:28

4:28 And all they in the synagogue, when they heard these things, were filled with wrath,

4:29 And rose up, and thrust him out of the city, and led et repleti sunt omnes in synagoga ira haec audientes 4:29

him unto the brow of the hill whereon their city was built, that they might cast him down headlong.

et surrexerunt et eiecerunt illum extra ciuitatem et duxerunt

Euangelium Secundum Lucam

illum usque ad supercilium montis supra quem ciuitas illorum erat aedificata ut praecipitarent eum 4:30 ipse autem 4:30 But he passing through the midst of them went his

way,

transiens per medium illorum ibat 4:31 et descendit in Ca- 4:31 And came down to Capernaum, a city of Galilee, and

taught them on the sabbath days.

pharnaum ciuitatem Galilaeae ibique docebat illos sabbatis 4:32 et stupebant in doctrina eius quia in potestate erat sermo 4:32 And they were astonished at his doctrine: for his

word was with power.

ipsius 4:33 et in synagoga erat homo habens daemonium 4:33 And in the synagogue there was a man, which had a

spirit of an unclean devil, and cried out with a loud voice,

inmundum et exclamauit uoce magna 4:34 dicens sine quid 4:34 Saying, Let us alone; what have we to do with thee,

thou Jesus of Nazareth? art thou come to destroy us? I know thee who thou art; the Holy One of God.

nobis et tibi Iesu Nazarene uenisti perdere nos scio te qui sis Sanctus Dei 4:35 et increpauit illi Iesus dicens obmu- 4:35 And Jesus rebuked him, saying, Hold thy peace, and

come out of him. And when the devil had thrown him in the midst, he came out of him, and hurt him not.

tesce et exi ab illo et cum proiecisset illum daemonium in medium exiit ab illo nihilque illum nocuit 4:36 et factus 4:36 And they were all amazed, and spake among them-

selves, saying, What a word is this! for with authority and power he commandeth the unclean spirits, and they come

est pauor in omnibus et conloquebantur ad inuicem dicen- out. tes quod est hoc uerbum quia in potestate et uirtute imperat inmundis spiritibus et exeunt 4:37 et diuulgabatur fama de 4:37 And the fame of him went out into every place of the

country round about.

illo in omnem locum regionis 4:38 surgens autem de sy- 4:38 And he arose out of the synagogue, and entered into Simon’s house. And Simon’s wife’s mother was taken with a great fever; and they besought him for her.

nagoga introiuit in domum Simonis socrus autem Simonis tenebatur magnis febribus et rogauerunt illum pro ea 4:39 4:39 And he stood over her, and rebuked the fever; and it

left her: and immediately she arose and ministered unto them.

et stans super illam imperauit febri et dimisit illam et continuo surgens ministrabat illis 4:40 cum sol autem occidisset 4:40 Now when the sun was setting, all they that had any

sick with divers diseases brought them unto him; and he laid his hands on every one of them, and healed them.

omnes qui habebant infirmos uariis languoribus ducebant illos ad eum at ille singulis manus inponens curabat eos 4:41 4:41 And devils also came out of many, crying out, and

saying, Thou art Christ the Son of God. And he rebuking them suffered them not to speak: for they knew that he

exiebant autem etiam daemonia a multis clamantia et dicen- was Christ. tia quia tu es Filius Dei et increpans non sinebat ea loqui quia sciebant ipsum esse Christum 4:42 facta autem die eg- 4:42 And when it was day, he departed and went into a

desert place: and the people sought him, and came unto him, and stayed him, that he should not depart from them.

ressus ibat in desertum locum et turbae requirebant eum et uenerunt usque ad ipsum et detinebant illum ne discederet ab eis 4:43 quibus ille ait quia et aliis ciuitatibus oportet 4:43 And he said unto them, I must preach the kingdom of

God to other cities also: for therefore am I sent.

me euangelizare regnum Dei quia ideo missus sum 4:44 et 4:44 And he preached in the synagogues of Galilee. erat praedicans in synagogis Galilaeae

5:1 factum est autem cum turbae inruerent in eum ut au- 5:1 And it came to pass, that, as the people pressed upon him to hear the word of God, he stood by the lake of Gennesaret,

dirent uerbum Dei et ipse stabat secus stagnum Gennesareth 5:2 et uidit duas naues stantes secus stagnum piscatores au- 5:2 And saw two ships standing by the lake: but the fisher-

men were gone out of them, and were washing their nets.

tem descenderant et lauabant retia 5:3 ascendens autem in 5:3 And he entered into one of the ships, which was Si-

mon’s, and prayed him that he would thrust out a little from the land. And he sat down, and taught the people out of the ship.

Biblia Sacra Vulgata

unam nauem quae erat Simonis rogauit eum a terra reducere

5:4 Now when he had left speaking, he said unto Simon, pusillum et sedens docebat de nauicula turbas 5:4 ut cessa-

Launch out into the deep, and let down your nets for a draught.

uit autem loqui dixit ad Simonem duc in altum et laxate retia

5:5 And Simon answering said unto him, Master, we have uestra in capturam 5:5 et respondens Simon dixit illi prae-

toiled all the night, and have taken nothing: nevertheless at thy word I will let down the net.

ceptor per totam noctem laborantes nihil cepimus in uerbo

5:6 And when they had this done, they inclosed a great autem tuo laxabo rete 5:6 et cum hoc fecissent concluse-

multitude of fishes: and their net brake.

runt piscium multitudinem copiosam rumpebatur autem rete

5:7 And they beckoned unto their partners, which were in eorum 5:7 et annuerunt sociis qui erant in alia naui ut ue-

the other ship, that they should come and help them. And they came, and filled both the ships, so that they began to

sink. nirent et adiuuarent eos et uenerunt et impleuerunt ambas

5:8 When Simon Peter saw it, he fell down at Jesus’ knees, nauiculas ita ut mergerentur 5:8 quod cum uideret Simon saying, Depart from me; for I am a sinful man, O Lord.

Petrus procidit ad genua Iesu dicens exi a me quia homo

5:9 For he was astonished, and all that were with him, at peccator sum Domine 5:9 stupor enim circumdederat eum the draught of the fishes which they had taken:

et omnes qui cum illo erant in captura piscium quam cepe-

5:10 And so was also James, and John, the sons of rant 5:10 similiter autem Iacobum et Iohannem filios Ze-

Zebedee, which were partners with Simon. And Jesus said unto Simon, Fear not; from henceforth thou shalt catch

men. bedaei qui erant socii Simonis et ait ad Simonem Iesus noli 5:11 And when they had brought their ships to land, they timere ex hoc iam homines eris capiens 5:11 et subductis

forsook all, and followed him.

5:12 And it came to pass, when he was in a certain city, ad terram nauibus relictis omnibus secuti sunt illum 5:12

behold a man full of leprosy: who seeing Jesus fell on his face, and besought him, saying, Lord, if thou wilt, thou

canst make me clean. et factum est cum esset in una ciuitatum et ecce uir plenus

lepra et uidens Iesum et procidens in faciem rogauit eum

5:13 And he put forth his hand, and touched him, saying, I dicens Domine si uis potes me mundare 5:13 et extendens will: be thou clean. And immediately the leprosy departed from him.

manum tetigit illum dicens uolo mundare et confestim lepra

5:14 And he charged him to tell no man: but go, and shew discessit ab illo 5:14 et ipse praecepit illi ut nemini diceret

thyself to the priest, and offer for thy cleansing, according as Moses commanded, for a testimony unto them.

sed uade ostende te sacerdoti et offer pro emundatione tua

5:15 But so much the more went there a fame abroad of sicut praecepit Moses in testimonium illis 5:15 perambula-

him: and great multitudes came together to hear, and to be healed by him of their infirmities.

bat autem magis sermo de illo et conueniebant turbae multae

5:16 And he withdrew himself into the wilderness, and ut audirent et curarentur ab infirmitatibus suis 5:16 ipse au-

prayed.

5:17 And it came to pass on a certain day, as he was teach- tem secedebat in deserto et orabat 5:17 et factum est in una

ing, that there were Pharisees and doctors of the law sitting by, which were come out of every town of Galilee, and Judaea, and Jerusalem: and the power of the Lord was present to heal them.

dierum et ipse sedebat docens et erant Pharisaei sedentes et legis doctores qui uenerant ex omni castello Galilaeae et Iudaeae et Hierusalem et uirtus erat Domini ad sanandum eos

5:18 And, behold, men brought in a bed a man which was 5:18 et ecce uiri portantes in lecto hominem qui erat paraly-

taken with a palsy: and they sought means to bring him in, and to lay him before him.

ticus et quaerebant eum inferre et ponere ante eum 5:19 et

5:19 And when they could not find by what way they might bring him in because of the multitude, they went upon the housetop, and let him down through the tiling with his couch into the midst before Jesus.

non inuenientes qua parte illum inferrent prae turba ascende-

Euangelium Secundum Lucam

runt supra tectum per tegulas submiserunt illum cum lecto in medium ante Iesum 5:20 quorum fidem ut uidit dixit homo 5:20 And when he saw their faith, he said unto him, Man,

thy sins are forgiven thee.

remittuntur tibi peccata tua 5:21 et coeperunt cogitare sc- 5:21 And the scribes and the Pharisees began to reason,

saying, Who is this which speaketh blasphemies? Who can forgive sins, but God alone?

ribae et Pharisaei dicentes quis est hic qui loquitur blasphemias quis potest dimittere peccata nisi solus Deus 5:22 ut 5:22 But when Jesus perceived their thoughts, he answer-

ing said unto them, What reason ye in your hearts?

cognouit autem Iesus cogitationes eorum respondens dixit ad illos quid cogitatis in cordibus uestris 5:23 quid est fa- 5:23 Whether is easier, to say, Thy sins be forgiven thee;

or to say, Rise up and walk?

cilius dicere dimittuntur tibi peccata an dicere surge et ambula 5:24 ut autem sciatis quia Filius hominis potestatem 5:24 But that ye may know that the Son of man hath power

upon earth to forgive sins, (he said unto the sick of the palsy,) I say unto thee, Arise, and take up thy couch, and

habet in terra dimittere peccata ait paralytico tibi dico surge go into thine house. tolle lectum tuum et uade in domum tuam 5:25 et confes- 5:25 And immediately he rose up before them, and took

up that whereon he lay, and departed to his own house, glorifying God.

tim surgens coram illis tulit in quo iacebat et abiit in domum suam magnificans Deum 5:26 et stupor adprehendit omnes 5:26 And they were all amazed, and they glorified God,

and were filled with fear, saying, We have seen strange things to day.

et magnificabant Deum et repleti sunt timore dicentes quia uidimus mirabilia hodie 5:27 et post haec exiit et uidit pu- 5:27 And after these things he went forth, and saw a pub-

lican, named Levi, sitting at the receipt of custom: and he said unto him, Follow me.

blicanum nomine Leui sedentem ad teloneum et ait illi sequere me 5:28 et relictis omnibus surgens secutus est eum 5:28 And he left all, rose up, and followed him. 5:29 et fecit ei conuiuium magnum Leui in domo sua et erat 5:29 And Levi made him a great feast in his own house:

and there was a great company of publicans and of others that sat down with them.

turba multa publicanorum et aliorum qui cum illis erant discumbentes 5:30 et murmurabant Pharisaei et scribae eorum 5:30 But their scribes and Pharisees murmured against his

disciples, saying, Why do ye eat and drink with publicans and sinners?

dicentes ad discipulos eius quare cum publicanis et peccatoribus manducatis et bibitis 5:31 et respondens Iesus dixit 5:31 And Jesus answering said unto them, They that are

whole need not a physician; but they that are sick.

ad illos non egent qui sani sunt medico sed qui male habent 5:32 non ueni uocare iustos sed peccatores in paenitentiam 5:32 I came not to call the righteous, but sinners to repen-

tance.

5:33 at illi dixerunt ad eum quare discipuli Iohannis ieiunant 5:33 And they said unto him, Why do the disciples of John

fast often, and make prayers, and likewise the disciples of the Pharisees; but thine eat and drink?

frequenter et obsecrationes faciunt similiter et Pharisaeorum tui autem edunt et bibunt 5:34 quibus ipse ait numquid po- 5:34 And he said unto them, Can ye make the children of

the bridechamber fast, while the bridegroom is with them?

testis filios sponsi dum cum illis est sponsus facere ieiunare 5:35 uenient autem dies et cum ablatus fuerit ab illis sponsus 5:35 But the days will come, when the bridegroom shall

be taken away from them, and then shall they fast in those days.

tunc ieiunabunt in illis diebus 5:36 dicebat autem et simili-

5:36 And he spake also a parable unto them; No man putteth a piece of a new garment upon an old; if otherwise, then both the new maketh a rent, and the piece that was taken out of the new agreeth not with the old.

tudinem ad illos quia nemo commissuram a uestimento nouo inmittit in uestimentum uetus alioquin et nouum rumpit et ueteri non conuenit commissura a nouo 5:37 et nemo mittit 5:37 And no man putteth new wine into old bottles; else

the new wine will burst the bottles, and be spilled, and the bottles shall perish.

Biblia Sacra Vulgata

uinum nouum in utres ueteres alioquin rumpet uinum nouum

5:38 But new wine must be put into new bottles; and both utres et ipsum effundetur et utres peribunt 5:38 sed uinum

are preserved.

nouum in utres nouos mittendum est et utraque conseruantur

5:39 No man also having drunk old wine straightway de- 5:39 et nemo bibens uetus statim uult nouum dicit enim uetus

sireth new: for he saith, The old is better.

melius est

6:1 And it came to pass on the second sabbath after the 6:1 factum est autem in sabbato secundoprimo cum trans-

first, that he went through the corn fields; and his disciples plucked the ears of corn, and did eat, rubbing them in their

hands. iret per sata uellebant discipuli eius spicas et manducabant 6:2 And certain of the Pharisees said unto them, Why do confricantes manibus 6:2 quidam autem Pharisaeorum di-

ye that which is not lawful to do on the sabbath days?

6:3 And Jesus answering them said, Have ye not read so cebant illis quid facitis quod non licet in sabbatis 6:3 et re-

much as this, what David did, when himself was an hungred, and they which were with him;

spondens Iesus ad eos dixit nec hoc legistis quod fecit Dauid

6:4 How he went into the house of God, and did take and cum esurisset ipse et qui cum eo erant 6:4 quomodo intrauit

eat the shewbread, and gave also to them that were with him; which it is not lawful to eat but for the priests alone?

in domum Dei et panes propositionis sumpsit et manducauit et dedit his qui cum ipso erant quos non licet manducare

6:5 And he said unto them, That the Son of man is Lord nisi tantum sacerdotibus 6:5 et dicebat illis quia dominus

also of the sabbath.

6:6 And it came to pass also on another sabbath, that he est Filius hominis etiam sabbati 6:6 factum est autem et in

entered into the synagogue and taught: and there was a man whose right hand was withered.

alio sabbato ut intraret in synagogam et doceret et erat ibi

6:7 And the scribes and Pharisees watched him, whether homo et manus eius dextra erat arida 6:7 obseruabant au-

he would heal on the sabbath day; that they might find an accusation against him.

tem scribae et Pharisaei si in sabbato curaret ut inuenirent

6:8 But he knew their thoughts, and said to the man which accusare illum 6:8 ipse uero sciebat cogitationes eorum et

had the withered hand, Rise up, and stand forth in the midst. And he arose and stood forth.

ait homini qui habebat manum aridam surge et sta in me-

6:9 Then said Jesus unto them, I will ask you one thing; Is dium et surgens stetit 6:9 ait autem ad illos Iesus interrogo

it lawful on the sabbath days to do good, or to do evil? to save life, or to destroy it?

uos si licet sabbato bene facere an male animam saluam fa-

6:10 And looking round about upon them all, he said unto cere an perdere 6:10 et circumspectis omnibus dixit homini

the man, Stretch forth thy hand. And he did so: and his hand was restored whole as the other.

extende manum tuam et extendit et restituta est manus eius

6:11 And they were filled with madness; and communed 6:11 ipsi autem repleti sunt insipientia et conloquebantur ad

one with another what they might do to Jesus.

6:12 And it came to pass in those days, that he went out inuicem quidnam facerent Iesu 6:12 factum est autem in

into a mountain to pray, and continued all night in prayer to God.

illis diebus exiit in montem orare et erat pernoctans in ora-

6:13 And when it was day, he called unto him his disci- tione Dei 6:13 et cum dies factus esset uocauit discipulos

ples: and of them he chose twelve, whom also he named apostles;

suos et elegit duodecim ex ipsis quos et apostolos nominauit

6:14 Simon, (whom he also named Peter,) and Andrew his 6:14 Simonem quem cognominauit Petrum et Andream frat-

brother, James and John, Philip and Bartholomew,

rem eius Iacobum et Iohannem Philippum et Bartholomeum

6:15 Matthew and Thomas, James the son of Alphaeus, 6:15 Mattheum et Thomam Iacobum Alphei et Simonem qui

and Simon called Zelotes, 6:16 And Judas the brother of James, and Judas Iscariot, uocatur Zelotes 6:16 Iudam Iacobi et Iudam Scarioth qui which also was the traitor.

Euangelium Secundum Lucam

fuit prodiior 6:17 et descendens cum illis stetit in loco cam- 6:17 And he came down with them, and stood in the plain,

and the company of his disciples, and a great multitude of people out of all Judaea and Jerusalem, and from the sea coast of Tyre and Sidon, which came to hear him, and to be healed of their diseases;

pestri et turba discipulorum eius et multitudo copiosa plebis ab omni Iudaea et Hierusalem et maritimae Tyri et Sidonis 6:18 qui uenerunt ut audirent eum et sanarentur a languori- 6:18 And they that were vexed with unclean spirits: and

they were healed.

bus suis et qui uexabantur ab spiritibus inmundis curabantur 6:19 et omnis turba quaerebant eum tangere quia uirtus de illo 6:19 And the whole multitude sought to touch him: for

there went virtue out of him, and healed them all.

exiebat et sanabat omnes 6:20 et ipse eleuatis oculis in dis- 6:20 And he lifted up his eyes on his disciples, and said,

Blessed be ye poor: for yours is the kingdom of God.

cipulos suos dicebat beati pauperes quia uestrum est regnum Dei 6:21 beati qui nunc esuritis quia saturabimini beati qui 6:21 Blessed are ye that hunger now: for ye shall be filled.

Blessed are ye that weep now: for ye shall laugh.

nunc fletis quia ridebitis 6:22 beati eritis cum uos oderint 6:22 Blessed are ye, when men shall hate you, and when

they shall separate you from their company, and shall reproach you, and cast out your name as evil, for the Son of

homines et cum separauerint uos et exprobrauerint et eie- man’s sake. cerint nomen uestrum tamquam malum propter Filium hominis 6:23 gaudete in illa die et exultate ecce enim merces 6:23 Rejoice ye in that day, and leap for joy: for, behold,

your reward is great in heaven: for in the like manner did their fathers unto the prophets.

uestra multa in caelo secundum haec enim faciebant prophetis patres eorum 6:24 uerumtamen uae uobis diuitibus quia 6:24 But woe unto you that are rich! for ye have received

your consolation.

habetis consolationem uestram 6:25 uae uobis qui saturati 6:25 Woe unto you that are full! for ye shall hunger. Woe

unto you that laugh now! for ye shall mourn and weep.

estis quia esurietis uae uobis qui ridetis nunc quia lugebitis et flebitis 6:26 uae cum bene uobis dixerint omnes homines 6:26 Woe unto you, when all men shall speak well of you!

for so did their fathers to the false prophets.

secundum haec faciebant prophetis patres eorum 6:27 sed 6:27 But I say unto you which hear, Love your enemies,

do good to them which hate you,

uobis dico qui auditis diligite inimicos uestros benefacite his qui uos oderunt 6:28 benedicite maledicentibus uobis orate 6:28 Bless them that curse you, and pray for them which

despitefully use you.

pro calumniantibus uos 6:29 ei qui te percutit in maxillam 6:29 And unto him that smiteth thee on the one cheek offer

also the other; and him that taketh away thy cloak forbid not to take thy coat also.

praebe et alteram et ab eo qui aufert tibi uestimentum etiam tunicam noli prohibere 6:30 omni autem petenti te tribue 6:30 Give to every man that asketh of thee; and of him that

taketh away thy goods ask them not again.

et qui aufert quae tua sunt ne repetas 6:31 et prout uultis 6:31 And as ye would that men should do to you, do ye

also to them likewise.

ut faciant uobis homines et uos facite illis similiter 6:32 et 6:32 For if ye love them which love you, what thank have

ye? for sinners also love those that love them.

si diligitis eos qui uos diligunt quae uobis est gratia nam et peccatores diligentes se diligunt 6:33 et si benefeceritis his 6:33 And if ye do good to them which do good to you,

what thank have ye? for sinners also do even the same.

qui uobis benefaciunt quae uobis est gratia siquidem et peccatores hoc faciunt 6:34 et si mutuum dederitis his a qui- 6:34 And if ye lend to them of whom ye hope to receive,

what thank have ye? for sinners also lend to sinners, to receive as much again.

bus speratis recipere quae gratia est uobis nam et peccatores peccatoribus fenerantur ut recipiant aequalia 6:35 uerumta- 6:35 But love ye your enemies, and do good, and lend,

hoping for nothing again; and your reward shall be great, and ye shall be the children of the Highest: for he is kind

men diligite inimicos uestros et benefacite et mutuum date unto the unthankful and to the evil.

Biblia Sacra Vulgata

nihil desperantes et erit merces uestra multa et eritis filii Al-

6:36 Be ye therefore merciful, as your Father also is mer- tissimi quia ipse benignus est super ingratos et malos 6:36

ciful.

estote ergo misericordes sicut et Pater uester misericors est

6:37 Judge not, and ye shall not be judged: condemn not, 6:37 nolite iudicare et non iudicabimini nolite condemnare

and ye shall not be condemned: forgive, and ye shall be forgiven:

et non condemnabimini dimittite et dimittemini 6:38 date

6:38 Give, and it shall be given unto you; good measure, pressed down, and shaken together, and running over, shall men give into your bosom. For with the same measure that ye mete withal it shall be measured to you again.

et dabitur uobis mensuram bonam confersam et coagitatam et supereffluentem dabunt in sinum uestrum eadem quippe

6:39 And he spake a parable unto them, Can the blind lead mensura qua mensi fueritis remetietur uobis 6:39 dicebat

the blind? shall they not both fall into the ditch?

autem illis et similitudinem numquid potest caecus caecum

6:40 The disciple is not above his master: but every one ducere nonne ambo in foueam cadent 6:40 non est discipu-

that is perfect shall be as his master.

lus super magistrum perfectus autem omnis erit sicut magis-

6:41 And why beholdest thou the mote that is in thy ter eius 6:41 quid autem uides festucam in oculo fratris tui

brother’s eye, but perceivest not the beam that is in thine own eye?

trabem autem quae in oculo tuo est non consideras 6:42 et

6:42 Either how canst thou say to thy brother, Brother, let me pull out the mote that is in thine eye, when thou thyself beholdest not the beam that is in thine own eye? Thou hypocrite, cast out first the beam out of thine own eye, and then shalt thou see clearly to pull out the mote that is in thy brother’s eye.

quomodo potes dicere fratri tuo frater sine eiciam festucam de oculo tuo ipse in oculo tuo trabem non uidens hypocrita eice primum trabem de oculo tuo et tunc perspicies ut educas

6:43 For a good tree bringeth not forth corrupt fruit; nei- festucam de oculo fratris tui 6:43 non est enim arbor bona ther doth a corrupt tree bring forth good fruit.

quae facit fructus malos neque arbor mala faciens fructum

6:44 For every tree is known by his own fruit. For of thorns bonum 6:44 unaquaeque enim arbor de fructu suo cognosmen do not gather figs, nor of a bramble bush gather they grapes.

citur neque enim de spinis colligunt ficus neque de rubo uin-

6:45 A good man out of the good treasure of his heart demiant uuam 6:45 bonus homo de bono thesauro cordis

bringeth forth that which is good; and an evil man out of the evil treasure of his heart bringeth forth that which is evil: for of the abundance of the heart his mouth speaketh. sui profert bonum et malus homo de malo profert malum ex

6:46 And why call ye me, Lord, Lord, and do not the abundantia enim cordis os loquitur 6:46 quid autem uoca-

things which I say?

6:47 Whosoever cometh to me, and heareth my sayings, tis me Domine Domine et non facitis quae dico 6:47 omnis

and doeth them, I will shew you to whom he is like:

qui uenit ad me et audit sermones meos et facit eos osten-

6:48 He is like a man which built an house, and digged dam uobis cui similis est 6:48 similis est homini aedificanti

deep, and laid the foundation on a rock: and when the flood arose, the stream beat vehemently upon that house,

and could not shake it: for it was founded upon a rock. domum qui fodit in altum et posuit fundamenta supra pet-

ram inundatione autem facta inlisum est flumen domui illi et

6:49 But he that heareth, and doeth not, is like a man that non potuit eam mouere fundata enim erat supra petram 6:49

without a foundation built an house upon the earth; against which the stream did beat vehemently, and immediately it

fell; and the ruin of that house was great. qui autem audiuit et non fecit similis est homini aedificanti domum suam supra terram sine fundamento in quam inlisus est fluuius et continuo concidit et facta est ruina domus illius magna

Euangelium Secundum Lucam

7:1 cum autem implesset omnia uerba sua in aures plebis 7:1 Now when he had ended all his sayings in the audience

of the people, he entered into Capernaum.

intrauit Capharnaum 7:2 centurionis autem cuiusdam ser- 7:2 And a certain centurion’s servant, who was dear unto

him, was sick, and ready to die.

uus male habens erat moriturus qui illi erat pretiosus 7:3 et 7:3 And when he heard of Jesus, he sent unto him the el-

ders of the Jews, beseeching him that he would come and heal his servant.

cum audisset de Iesu misit ad eum seniores Iudaeorum rogans eum ut ueniret et saluaret seruum eius 7:4 at illi cum 7:4 And when they came to Jesus, they besought him in-

stantly, saying, That he was worthy for whom he should do this:

uenissent ad Iesum rogabant eum sollicite dicentes ei quia dignus est ut hoc illi praestes 7:5 diligit enim gentem no- 7:5 For he loveth our nation, and he hath built us a syna-

gogue.

stram et synagogam ipse aedificauit nobis 7:6 Iesus autem 7:6 Then Jesus went with them. And when he was now

not far from the house, the centurion sent friends to him, saying unto him, Lord, trouble not thyself: for I am not

ibat cum illis et cum iam non longe esset a domo misit ad worthy that thou shouldest enter under my roof: eum centurio amicos dicens Domine noli uexari non enim dignus sum ut sub tectum meum intres 7:7 propter quod et 7:7 Wherefore neither thought I myself worthy to come

unto thee: but say in a word, and my servant shall be healed.

me ipsum non sum dignum arbitratus ut uenirem ad te sed

dic uerbo et sanabitur puer meus 7:8 nam et ego homo sum 7:8 For I also am a man set under authority, having under me soldiers, and I say unto one, Go, and he goeth; and to another, Come, and he cometh; and to my servant, Do this,

sub potestate constitutus habens sub me milites et dico huic and he doeth it. uade et uadit et alio ueni et uenit et seruo meo fac hoc et facit 7:9 quo audito Iesus miratus est et conuersus sequen- 7:9 When Jesus heard these things, he marvelled at him,

and turned him about, and said unto the people that followed him, I say unto you, I have not found so great faith,

tibus se turbis dixit amen dico uobis nec in Israhel tantam no, not in Israel.

fidem inueni 7:10 et reuersi qui missi fuerant domum inue- 7:10 And they that were sent, returning to the house, found

the servant whole that had been sick.

nerunt seruum qui languerat sanum 7:11 et factum est dein- 7:11 And it came to pass the day after, that he went into a

city called Nain; and many of his disciples went with him, and much people.

ceps ibat in ciuitatem quae uocatur Naim et ibant cum illo discipuli eius et turba copiosa 7:12 cum autem adpropinqua- 7:12 Now when he came nigh to the gate of the city, be-

hold, there was a dead man carried out, the only son of his mother, and she was a widow: and much people of the city

ret portae ciuitatis et ecce defunctus efferebatur filius unicus was with her. matri suae et haec uidua erat et turba ciuitatis multa cum illa 7:13 quam cum uidisset Dominus misericordia motus 7:13 And when the Lord saw her, he had compassion on

her, and said unto her, Weep not.

super ea dixit illi noli flere 7:14 et accessit et tetigit locu- 7:14 And he came and touched the bier: and they that bare

him stood still. And he said, Young man, I say unto thee, Arise.

lum hii autem qui portabant steterunt et ait adulescens tibi dico surge 7:15 et resedit qui erat mortuus et coepit loqui 7:15 And he that was dead sat up, and began to speak. And

he delivered him to his mother.

et dedit illum matri suae 7:16 accepit autem omnes timor et 7:16 And there came a fear on all: and they glorified God,

saying, That a great prophet is risen up among us; and, That God hath visited his people.

magnificabant Deum dicentes quia propheta magnus surrexit in nobis et quia Deus uisitauit plebem suam 7:17 et exiit hic 7:17 And this rumour of him went forth throughout all

Judaea, and throughout all the region round about.

sermo in uniuersam Iudaeam de eo et omnem circa regionem 7:18 et nuntiauerunt Iohanni discipuli eius de omni- 7:18 And the disciples of John shewed him of all these

things.

bus his 7:19 et conuocauit duos de discipulis suis Iohannes 7:19 And John calling unto him two of his disciples sent

them to Jesus, saying, Art thou he that should come? or look we for another?

Biblia Sacra Vulgata

et misit ad Dominum dicens tu es qui uenturus es an alium

7:20 When the men were come unto him, they said, John expectamus 7:20 cum autem uenissent ad eum uiri dixerunt

Baptist hath sent us unto thee, saying, Art thou he that should come? or look we for another?

Iohannes Baptista misit nos ad te dicens tu es qui uenturus es

7:21 And in that same hour he cured many of their infir- an alium expectamus 7:21 in ipsa autem hora curauit mul-

mities and plagues, and of evil spirits; and unto many that were blind he gave sight.

tos a languoribus et plagis et spiritibus malis et caecis mul-

7:22 Then Jesus answering said unto them, Go your way, tis donauit uisum 7:22 et respondens dixit illis euntes nun-

and tell John what things ye have seen and heard; how that the blind see, the lame walk, the lepers are cleansed, the deaf hear, the dead are raised, to the poor the gospel is preached.

tiate Iohanni quae uidistis et audistis quia caeci uident claudi ambulant leprosi mundantur surdi audiunt mortui resurgunt

7:23 And blessed is he, whosoever shall not be offended pauperes euangelizantur 7:23 et beatus est quicumque non

in me.

7:24 And when the messengers of John were departed, he fuerit scandalizatus in me 7:24 et cum discessissent nun-

began to speak unto the people concerning John, What went ye out into the wilderness for to see? A reed shaken

with the wind? tii Iohannis coepit dicere de Iohanne ad turbas quid existis 7:25 But what went ye out for to see? A man clothed in in desertum uidere harundinem uento moueri 7:25 sed quid

soft raiment? Behold, they which are gorgeously apparelled, and live delicately, are in kings’ courts.

existis uidere hominem mollibus uestimentis indutum ecce qui in ueste pretiosa sunt et deliciis in domibus regum sunt

7:26 But what went ye out for to see? A prophet? Yea, I 7:26 sed quid existis uidere prophetam utique dico uobis et

say unto you, and much more than a prophet.

7:27 This is he, of whom it is written, Behold, I send my plus quam prophetam 7:27 hic est de quo scriptum est ecce

messenger before thy face, which shall prepare thy way before thee.

mitto angelum meum ante faciem tuam qui praeparabit uiam

7:28 For I say unto you, Among those that are born of tuam ante te 7:28 dico enim uobis maior inter natos mulie-

women there is not a greater prophet than John the Baptist: but he that is least in the kingdom of God is greater than

he. rum propheta Iohanne Baptista nemo est qui autem minor

7:29 And all the people that heard him, and the publicans, est in regno Dei maior est illo 7:29 et omnis populus aujustified God, being baptized with the baptism of John.

diens et publicani iustificauerunt Deum baptizati baptismo

7:30 But the Pharisees and lawyers rejected the counsel of Iohannis 7:30 Pharisaei autem et legis periti consilium Dei

God against themselves, being not baptized of him.

7:31 And the Lord said, Whereunto then shall I liken the spreuerunt in semet ipsos non baptizati ab eo 7:31 cui ergo

men of this generation? and to what are they like?

similes dicam homines generationis huius et cui similes sunt

7:32 They are like unto children sitting in the marketplace, 7:32 similes sunt pueris sedentibus in foro et loquentibus ad

and calling one to another, and saying, We have piped unto you, and ye have not danced; we have mourned to you, and

ye have not wept. inuicem et dicentibus cantauimus uobis tibiis et non salta-7:33 For John the Baptist came neither eating bread nor stis lamentauimus et non plorastis 7:33 uenit enim Iohan-

drinking wine; and ye say, He hath a devil.

nes Baptista neque manducans panem neque bibens uinum

7:34 The Son of man is come eating and drinking; and ye et dicitis daemonium habet 7:34 uenit Filius hominis man-

say, Behold a gluttonous man, and a winebibber, a friend of publicans and sinners!

ducans et bibens et dicitis ecce homo deuorator et bibens ui-

7:35 But wisdom is justified of all her children. num amicus publicanorum et peccatorum 7:35 et iustificata 7:36 And one of the Pharisees desired him that he would est sapientia ab omnibus filiis suis 7:36 rogabat autem illum

eat with him. And he went into the Pharisee’s house, and sat down to meat.

quidam de Pharisaeis ut manducaret cum illo et ingressus

Euangelium Secundum Lucam

domum Pharisaei discubuit 7:37 et ecce mulier quae erat in 7:37 And, behold, a woman in the city, which was a sin-

ner, when she knew that Jesus sat at meat in the Pharisee’s house, brought an alabaster box of ointment,

ciuitate peccatrix ut cognouit quod accubuit in domo Pharisaei adtulit alabastrum unguenti 7:38 et stans retro secus 7:38 And stood at his feet behind him weeping, and began

to wash his feet with tears, and did wipe them with the hairs of her head, and kissed his feet, and anointed them

pedes eius lacrimis coepit rigare pedes eius et capillis ca- with the ointment. pitis sui tergebat et osculabatur pedes eius et unguento unguebat 7:39 uidens autem Pharisaeus qui uocauerat eum ait 7:39 Now when the Pharisee which had bidden him saw

it, he spake within himself, saying, This man, if he were a prophet, would have known who and what manner of

intra se dicens hic si esset propheta sciret utique quae et qua- woman this is that toucheth him: for she is a sinner. lis mulier quae tangit eum quia peccatrix est 7:40 et res- 7:40 And Jesus answering said unto him, Simon, I have

somewhat to say unto thee. And he saith, Master, say on.

pondens Iesus dixit ad illum Simon habeo tibi aliquid dicere at ille ait magister dic 7:41 duo debitores erant cuidam 7:41 There was a certain creditor which had two debtors:

the one owed five hundred pence, and the other fifty.

feneratori unus debebat denarios quingentos alius quinquaginta 7:42 non habentibus illis unde redderent donauit ut- 7:42 And when they had nothing to pay, he frankly forgave

them both. Tell me therefore, which of them will love him most?

risque quis ergo eum plus diliget 7:43 respondens Simon

7:43 Simon answered and said, I suppose that he, to whom he forgave most. And he said unto him, Thou hast rightly judged.

dixit aestimo quia is cui plus donauit at ille dixit ei recte iudicasti 7:44 et conuersus ad mulierem dixit Simoni ui- 7:44 And he turned to the woman, and said unto Simon,

Seest thou this woman? I entered into thine house, thou gavest me no water for my feet: but she hath washed my

des hanc mulierem intraui in domum tuam aquam pedibus feet with tears, and wiped them with the hairs of her head.

meis non dedisti haec autem lacrimis rigauit pedes meos et capillis suis tersit 7:45 osculum mihi non dedisti haec au- 7:45 Thou gavest me no kiss: but this woman since the

time I came in hath not ceased to kiss my feet.

tem ex quo intrauit non cessauit osculari pedes meos 7:46 7:46 My head with oil thou didst not anoint: but this

woman hath anointed my feet with ointment.

oleo caput meum non unxisti haec autem unguento unxit pedes meos 7:47 propter quod dico tibi remittentur ei peccata 7:47 Wherefore I say unto thee, Her sins, which are many,

are forgiven; for she loved much: but to whom little is forgiven, the same loveth little.

multa quoniam dilexit multum cui autem minus dimittitur minus diligit 7:48 dixit autem ad illam remittuntur tibi pec- 7:48 And he said unto her, Thy sins are forgiven. cata 7:49 et coeperunt qui simul accumbebant dicere intra 7:49 And they that sat at meat with him began to say

within themselves, Who is this that forgiveth sins also?

se quis est hic qui etiam peccata dimittit 7:50 dixit autem 7:50 And he said to the woman, Thy faith hath saved thee;

go in peace.

ad mulierem fides tua te saluam fecit uade in pace

8:1 et factum est deinceps et ipse iter faciebat per ciuita- 8:1 And it came to pass afterward, that he went through-

out every city and village, preaching and shewing the glad tidings of the kingdom of God: and the twelve were with

tem et castellum praedicans et euangelizans regnum Dei et him, duodecim cum illo 8:2 et mulieres aliquae quae erant cura- 8:2 And certain women, which had been healed of evil

spirits and infirmities, Mary called Magdalene, out of whom went seven devils,

tae ab spiritibus malignis et infirmitatibus Maria quae uocatur Magdalene de qua daemonia septem exierant 8:3 et Io- 8:3 And Joanna the wife of Chuza Herod’s steward, and

Susanna, and many others, which ministered unto him of their substance.

hanna uxor Chuza procuratoris Herodis et Susanna et aliae multae quae ministrabant eis de facultatibus suis 8:4 cum 8:4 And when much people were gathered together, and

were come to him out of every city, he spake by a parable:

Biblia Sacra Vulgata

autem turba plurima conueniret et de ciuitatibus properarent

8:5 A sower went out to sow his seed: and as he sowed, ad eum dixit per similitudinem 8:5 exiit qui seminat semi-

some fell by the way side; and it was trodden down, and the fowls of the air devoured it.

nare semen suum et dum seminat aliud cecidit secus uiam et

8:6 And some fell upon a rock; and as soon as it was conculcatum est et uolucres caeli comederunt illud 8:6 et

sprung up, it withered away, because it lacked moisture.

aliud cecidit supra petram et natum aruit quia non habebat

8:7 And some fell among thorns; and the thorns sprang up humorem 8:7 et aliud cecidit inter spinas et simul exortae

with it, and choked it.

8:8 And other fell on good ground, and sprang up, and spinae suffocauerunt illud 8:8 et aliud cecidit in terram bo-

bare fruit an hundredfold. And when he had said these things, he cried, He that hath ears to hear, let him hear.

nam et ortum fecit fructum centuplum haec dicens clama-

8:9 And his disciples asked him, saying, What might this bat qui habet aures audiendi audiat 8:9 interrogabant au-

parable be?

8:10 And he said, Unto you it is given to know the mys- tem eum discipuli eius quae esset haec parabola 8:10 qui-

teries of the kingdom of God: but to others in parables; that seeing they might not see, and hearing they might not

understand. bus ipse dixit uobis datum est nosse mysterium regni Dei

ceteris autem in parabolis ut uidentes non uideant et audien-

8:11 Now the parable is this: The seed is the word of God. tes non intellegant 8:11 est autem haec parabola semen est

8:12 Those by the way side are they that hear; then cometh uerbum Dei 8:12 qui autem secus uiam sunt qui audiunt dethe devil, and taketh away the word out of their hearts, lest they should believe and be saved.

inde uenit diabolus et tollit uerbum de corde eorum ne cre-

8:13 They on the rock are they, which, when they hear, dentes salui fiant 8:13 nam qui supra petram qui cum audie-

receive the word with joy; and these have no root, which for a while believe, and in time of temptation fall away.

rint cum gaudio suscipiunt uerbum et hii radices non habent qui ad tempus credunt et in tempore temptationis recedunt

8:14 And that which fell among thorns are they, which, 8:14 quod autem in spinis cecidit hii sunt qui audierunt et a

when they have heard, go forth, and are choked with cares and riches and pleasures of this life, and bring no fruit to

perfection. sollicitudinibus et diuitiis et uoluptatibus uitae euntes suf-8:15 But that on the good ground are they, which in an focantur et non referunt fructum 8:15 quod autem in bo-

honest and good heart, having heard the word, keep it, and bring forth fruit with patience.

nam terram hii sunt qui in corde bono et optimo audientes

8:16 No man, when he hath lighted a candle, covereth it uerbum retinent et fructum adferunt in patientia 8:16 nemo

with a vessel, or putteth it under a bed; but setteth it on a candlestick, that they which enter in may see the light.

autem lucernam accendens operit eam uaso aut subtus lectum ponit sed supra candelabrum ponit ut intrantes uideant

8:17 For nothing is secret, that shall not be made manifest; lumen 8:17 non enim est occultum quod non manifestetur neither any thing hid, that shall not be known and come abroad.

nec absconditum quod non cognoscatur et in palam ueniat

8:18 Take heed therefore how ye hear: for whosoever hath, 8:18 uidete ergo quomodo auditis qui enim habet dabitur illi

to him shall be given; and whosoever hath not, from him shall be taken even that which he seemeth to have.

et quicumque non habet etiam quod putat se habere aufe-

8:19 Then came to him his mother and his brethren, and retur ab illo 8:19 uenerunt autem ad illum mater et fratres

could not come at him for the press.

8:20 And it was told him by certain which said, Thy eius et non poterant adire ad eum prae turba 8:20 et nun-

mother and thy brethren stand without, desiring to see thee.

tiatum est illi mater tua et fratres tui stant foris uolentes te

8:21 And he answered and said unto them, My mother and uidere 8:21 qui respondens dixit ad eos mater mea et frat-

my brethren are these which hear the word of God, and do it.

Euangelium Secundum Lucam

res mei hii sunt qui uerbum Dei audiunt et faciunt 8:22 fac- 8:22 Now it came to pass on a certain day, that he went into

a ship with his disciples: and he said unto them, Let us go over unto the other side of the lake. And they launched

tum est autem in una dierum et ipse ascendit in nauiculam forth. et discipuli eius et ait ad illos transfretemus trans stagnum et ascenderunt 8:23 nauigantibus autem illis obdormiit et 8:23 But as they sailed he fell asleep: and there came down

a storm of wind on the lake; and they were filled with water, and were in jeopardy.

descendit procella uenti in stagnum et conplebantur et periclitabantur 8:24 accedentes autem suscitauerunt eum dicen- 8:24 And they came to him, and awoke him, saying, Mas-

ter, master, we perish. Then he arose, and rebuked the wind and the raging of the water: and they ceased, and

tes praeceptor perimus at ille surgens increpauit uentum et there was a calm. tempestatem aquae et cessauit et facta est tranquillitas 8:25 8:25 And he said unto them, Where is your faith? And

they being afraid wondered, saying one to another, What manner of man is this! for he commandeth even the winds

dixit autem illis ubi est fides uestra qui timentes mirati sunt and water, and they obey him. dicentes ad inuicem quis putas hic est quia et uentis imperat et mari et oboediunt ei 8:26 enauigauerunt autem ad regio- 8:26 And they arrived at the country of the Gadarenes,

which is over against Galilee.

nem Gerasenorum quae est contra Galilaeam 8:27 et cum 8:27 And when he went forth to land, there met him out

of the city a certain man, which had devils long time, and ware no clothes, neither abode in any house, but in the

egressus esset ad terram occurrit illi uir quidam qui habebat tombs. daemonium iam temporibus multis et uestimento non indue-

batur neque in domo manebat sed in monumentis 8:28 is 8:28 When he saw Jesus, he cried out, and fell down before him, and with a loud voice said, What have I to do with thee, Jesus, thou Son of God most high? I beseech thee,

ut uidit Iesum procidit ante illum et exclamans uoce magna torment me not. dixit quid mihi et tibi est Iesu Fili Dei altissimi obsecro te ne me torqueas 8:29 praecipiebat enim spiritui inmundo ut 8:29 (For he had commanded the unclean spirit to come

out of the man. For oftentimes it had caught him: and he was kept bound with chains and in fetters; and he brake the

exiret ab homine multis enim temporibus arripiebat illum et bands, and was driven of the devil into the wilderness.) uinciebatur catenis et conpedibus custoditus et ruptis uinculis agebatur a daemonio in deserta 8:30 interrogauit autem 8:30 And Jesus asked him, saying, What is thy name? And

he said, Legion: because many devils were entered into him.

illum Iesus dicens quod tibi nomen est at ille dixit Legio quia intrauerunt daemonia multa in eum 8:31 et rogabant 8:31 And they besought him that he would not command

them to go out into the deep.

illum ne imperaret illis ut in abyssum irent 8:32 erat autem 8:32 And there was there an herd of many swine feeding

on the mountain: and they besought him that he would suffer them to enter into them. And he suffered them.

ibi grex porcorum multorum pascentium in monte et rogabant eum ut permitteret eos in illos ingredi et permisit illos 8:33 exierunt ergo daemonia ab homine et intrauerunt 8:33 Then went the devils out of the man, and entered into

the swine: and the herd ran violently down a steep place into the lake, and were choked.

in porcos et impetu abiit grex per praeceps in stagnum et suffocatus est 8:34 quod ut uiderunt factum qui pascebant 8:34 When they that fed them saw what was done, they

fled, and went and told it in the city and in the country.

fugerunt et nuntiauerunt in ciuitatem et in uillas 8:35 exie- 8:35 Then they went out to see what was done; and came

to Jesus, and found the man, out of whom the devils were departed, sitting at the feet of Jesus, clothed, and in his

runt autem uidere quod factum est et uenerunt ad Iesum et right mind: and they were afraid. inuenerunt hominem sedentem a quo daemonia exierant uestitum ac sana mente ad pedes eius et timuerunt 8:36 nuntia- 8:36 They also which saw it told them by what means he

that was possessed of the devils was healed.

Biblia Sacra Vulgata

uerunt autem illis et qui uiderant quomodo sanus factus esset

8:37 Then the whole multitude of the country of the a Legione 8:37 et rogauerunt illum omnis multitudo regio-

Gadarenes round about besought him to depart from them; for they were taken with great fear: and he went up into

the ship, and returned back again. nis Gerasenorum ut discederet ab ipsis quia timore magno 8:38 Now the man out of whom the devils were departed tenebantur ipse autem ascendens nauem reuersus est 8:38

besought him that he might be with him: but Jesus sent him away, saying,

et rogabat illum uir a quo daemonia exierant ut cum eo es-

8:39 Return to thine own house, and shew how great things set dimisit autem eum Iesus dicens 8:39 redi domum tuam God hath done unto thee. And he went his way, and published throughout the whole city how great things Jesus

had done unto him. et narra quanta tibi fecit Deus et abiit per uniuersam ciui-8:40 And it came to pass, that, when Jesus was returned, tatem praedicans quanta illi fecisset Iesus 8:40 factum est

the people gladly received him: for they were all waiting for him.

autem cum rediiset Iesus excepit illum turba erant enim om-

8:41 And, behold, there came a man named Jairus, and he nes expectantes eum 8:41 et ecce uenit uir cui nomen Iairus

was a ruler of the synagogue: and he fell down at Jesus’ feet, and besought him that he would come into his house:

et ipse princeps synagogae erat et cecidit ad pedes Iesu ro-

8:42 For he had one only daughter, about twelve years gans eum ut intraret in domum eius 8:42 quia filia unica

of age, and she lay a dying. But as he went the people thronged him.

erat illi fere annorum duodecim et haec moriebatur et conti-

8:43 And a woman having an issue of blood twelve years, git dum iret a turbis conprimebatur 8:43 et mulier quaedam

which had spent all her living upon physicians, neither could be healed of any,

erat in fluxu sanguinis ab annis duodecim quae in medicos erogauerat omnem substantiam suam nec ab ullo potuit cu-

8:44 Came behind him, and touched the border of his gar- rari 8:44 accessit retro et tetigit fimbriam uestimenti eius et

ment: and immediately her issue of blood stanched.

8:45 And Jesus said, Who touched me? When all denied, confestim stetit fluxus sanguinis eius 8:45 et ait Iesus quis

Peter and they that were with him said, Master, the multitude throng thee and press thee, and sayest thou, Who

touched me? est qui me tetigit negantibus autem omnibus dixit Petrus et

qui cum illo erant praeceptor turbae te conprimunt et adfli-

8:46 And Jesus said, Somebody hath touched me: for I gunt et dicis quis me tetigit 8:46 et dixit Iesus tetigit me ali-

perceive that virtue is gone out of me.

8:47 And when the woman saw that she was not hid, she quis nam ego noui uirtutem de me exisse 8:47 uidens autem

came trembling, and falling down before him, she declared unto him before all the people for what cause she had

touched him, and how she was healed immediately. mulier quia non latuit tremens uenit et procidit ante pedes il-

lius et ob quam causam tetigerit eum indicauit coram omni

8:48 And he said unto her, Daughter, be of good comfort: populo et quemadmodum confestim sanata sit 8:48 at ipse

thy faith hath made thee whole; go in peace.

8:49 While he yet spake, there cometh one from the ruler dixit illi filia fides tua te saluam fecit uade in pace 8:49 ad-

of the synagogue’s house, saying to him, Thy daughter is dead; trouble not the Master.

huc illo loquente uenit a principe synagogae dicens ei quia

8:50 But when Jesus heard it, he answered him, saying, mortua est filia tua noli uexare illum 8:50 Iesus autem au-

Fear not: believe only, and she shall be made whole.

dito hoc uerbo respondit patri puellae noli timere crede tan-

8:51 And when he came into the house, he suffered no man tum et salua erit 8:51 et cum uenisset domum non permisit

to go in, save Peter, and James, and John, and the father and the mother of the maiden.

intrare secum quemquam nisi Petrum et Iohannem et Iaco-

8:52 And all wept, and bewailed her: but he said, Weep bum et patrem et matrem puellae 8:52 flebant autem omnes

not; she is not dead, but sleepeth.

et plangebant illam at ille dixit nolite flere non est mortua sed

Euangelium Secundum Lucam

dormit 8:53 et deridebant eum scientes quia mortua esset 8:53 And they laughed him to scorn, knowing that she was

dead.

8:54 ipse autem tenens manum eius clamauit dicens puella 8:54 And he put them all out, and took her by the hand,

and called, saying, Maid, arise.

surge 8:55 et reuersus est spiritus eius et surrexit continuo 8:55 And her spirit came again, and she arose straightway:

and he commanded to give her meat.

et iussit illi dari manducare 8:56 et stupuerunt parentes eius 8:56 And her parents were astonished: but he charged

them that they should tell no man what was done.

quibus praecepit ne alicui dicerent quod factum erat

9:1 conuocatis autem duodecim apostolis dedit illis uir- 9:1 Then he called his twelve disciples together, and gave

them power and authority over all devils, and to cure diseases.

tutem et potestatem super omnia daemonia et ut languores curarent 9:2 et misit illos praedicare regnum Dei et sanare 9:2 And he sent them to preach the kingdom of God, and

to heal the sick.

infirmos 9:3 et ait ad illos nihil tuleritis in uia neque uir- 9:3 And he said unto them, Take nothing for your journey,

neither staves, nor scrip, neither bread, neither money; neither have two coats apiece.

gam neque peram neque panem neque pecuniam neque duas tunicas habeatis 9:4 et in quamcumque domum intraueritis 9:4 And whatsoever house ye enter into, there abide, and

thence depart.

ibi manete et inde ne exeatis 9:5 et quicumque non recepe- 9:5 And whosoever will not receive you, when ye go out

of that city, shake off the very dust from your feet for a testimony against them.

rint uos exeuntes de ciuitate illa etiam puluerem pedum uestrorum excutite in testimonium supra illos 9:6 egressi au- 9:6 And they departed, and went through the towns,

preaching the gospel, and healing every where.

tem circumibant per castella euangelizantes et curantes ubique 9:7 audiuit autem Herodes tetrarcha omnia quae fie- 9:7 Now Herod the tetrarch heard of all that was done by

him: and he was perplexed, because that it was said of some, that John was risen from the dead;

bant ab eo et haesitabat eo quod diceretur 9:8 a quibusdam

9:8 And of some, that Elias had appeared; and of others, that one of the old prophets was risen again.

quia Iohannes surrexit a mortuis a quibusdam uero quia Helias apparuit ab aliis autem quia propheta unus de antiquis surrexit 9:9 et ait Herodes Iohannem ego decollaui quis au- 9:9 And Herod said, John have I beheaded: but who is this,

of whom I hear such things? And he desired to see him.

tem est iste de quo audio ego talia et quaerebat uidere eum 9:10 et reuersi apostoli narrauerunt illi quaecumque fecerunt 9:10 And the apostles, when they were returned, told him

all that they had done. And he took them, and went aside privately into a desert place belonging to the city called

et adsumptis illis secessit seorsum in locum desertum qui est Bethsaida. Bethsaida 9:11 quod cum cognouissent turbae secutae sunt 9:11 And the people, when they knew it, followed him:

and he received them, and spake unto them of the kingdom of God, and healed them that had need of healing.

illum et excepit illos et loquebatur illis de regno Dei et eos qui cura indigebant sanabat 9:12 dies autem coeperat de- 9:12 And when the day began to wear away, then came the

twelve, and said unto him, Send the multitude away, that they may go into the towns and country round about, and

clinare et accedentes duodecim dixerunt illi dimitte turbas lodge, and get victuals: for we are here in a desert place. ut euntes in castella uillasque quae circa sunt deuertant et inueniant escas quia hic in loco deserto sumus 9:13 ait au- 9:13 But he said unto them, Give ye them to eat. And

they said, We have no more but five loaves and two fishes; except we should go and buy meat for all this people.

tem ad illos uos date illis manducare at illi dixerunt non sunt nobis plus quam quinque panes et duo pisces nisi forte nos eamus et emamus in omnem hanc turbam escas 9:14 erant 9:14 For they were about five thousand men. And he said

to his disciples, Make them sit down by fifties in a company.

autem fere uiri quinque milia ait autem ad discipulos suos

Biblia Sacra Vulgata

9:15 And they did so, and made them all sit down. facite illos discumbere per conuiuia quinquagenos 9:15 et 9:16 Then he took the five loaves and the two fishes, and ita fecerunt et discumbere fecerunt omnes 9:16 acceptis au-

looking up to heaven, he blessed them, and brake, and gave to the disciples to set before the multitude.

tem quinque panibus et duobus piscibus respexit in caelum et benedixit illis et fregit et distribuit discipulis suis ut po-

9:17 And they did eat, and were all filled: and there was nerent ante turbas 9:17 et manducauerunt omnes et satu-

taken up of fragments that remained to them twelve baskets.

rati sunt et sublatum est quod superfuit illis fragmentorum

9:18 And it came to pass, as he was alone praying, his dis- cofini duodecim 9:18 et factum est cum solus esset orans ciples were with him: and he asked them, saying, Whom say the people that I am?

erant cum illo et discipuli et interrogauit illos dicens quem

9:19 They answering said, John the Baptist; but some say, me dicunt esse turbae 9:19 at illi responderunt et dixerunt

Elias; and others say, that one of the old prophets is risen again.

Iohannem Baptistam alii autem Heliam alii quia propheta

9:20 He said unto them, But whom say ye that I am? Peter unus de prioribus surrexit 9:20 dixit autem illis uos autem

answering said, The Christ of God.

quem me esse dicitis respondens Simon Petrus dixit Chris-

9:21 And he straitly charged them, and commanded them tum Dei 9:21 at ille increpans illos praecepit ne cui dicerent

to tell no man that thing;

9:22 Saying, The Son of man must suffer many things, and hoc 9:22 dicens quia oportet Filium hominis multa pati et

be rejected of the elders and chief priests and scribes, and be slain, and be raised the third day.

reprobari a senioribus et principibus sacerdotum et scribis et

9:23 And he said to them all, If any man will come after occidi et tertia die resurgere 9:23 dicebat autem ad omnes

me, let him deny himself, and take up his cross daily, and follow me.

si quis uult post me uenire abneget se ipsum et tollat crucem

9:24 For whosoever will save his life shall lose it: but suam cotidie et sequatur me 9:24 qui enim uoluerit animam

whosoever will lose his life for my sake, the same shall save it.

suam saluam facere perdet illam nam qui perdiderit animam

9:25 For what is a man advantaged, if he gain the whole suam propter me saluam faciet illam 9:25 quid enim pro-

world, and lose himself, or be cast away?

ficit homo si lucretur uniuersum mundum se autem ipsum

9:26 For whosoever shall be ashamed of me and of my perdat et detrimentum sui faciat 9:26 nam qui me erubuerit

words, of him shall the Son of man be ashamed, when he shall come in his own glory, and in his Father’s, and of the

holy angels. et meos sermones hunc Filius hominis erubescet cum uene-9:27 But I tell you of a truth, there be some standing here, rit in maiestate sua et Patris et sanctorum angelorum 9:27

which shall not taste of death, till they see the kingdom of God.

dico autem uobis uere sunt aliqui hic stantes qui non gus-

9:28 And it came to pass about an eight days after these tabunt mortem donec uideant regnum Dei 9:28 factum est

sayings, he took Peter and John and James, and went up into a mountain to pray.

autem post haec uerba fere dies octo et adsumpsit Petrum et

9:29 And as he prayed, the fashion of his countenance was Iohannem et Iacobum et ascendit in montem ut oraret 9:29 altered, and his raiment was white and glistering.

et factum est dum oraret species uultus eius altera et uesti-

9:30 And, behold, there talked with him two men, which tus eius albus refulgens 9:30 et ecce duo uiri loquebantur

were Moses and Elias:

9:31 Who appeared in glory, and spake of his decease cum illo erant autem Moses et Helias 9:31 uisi in maiestate

which he should accomplish at Jerusalem.

et dicebant excessum eius quem conpleturus erat in Hieru-

9:32 But Peter and they that were with him were heavy salem 9:32 Petrus uero et qui cum illo grauati erant somno

with sleep: and when they were awake, they saw his glory, and the two men that stood with him.

Euangelium Secundum Lucam

et euigilantes uiderunt maiestatem eius et duos uiros qui stabant cum illo 9:33 et factum est cum discederent ab illo ait 9:33 And it came to pass, as they departed from him, Peter

said unto Jesus, Master, it is good for us to be here: and let us make three tabernacles; one for thee, and one for

Petrus ad Iesum praeceptor bonum est nos hic esse et fa- Moses, and one for Elias: not knowing what he said. ciamus tria tabernacula unum tibi et unum Mosi et unum Heliae nesciens quid diceret 9:34 haec autem illo loquente 9:34 While he thus spake, there came a cloud, and over-

shadowed them: and they feared as they entered into the cloud.

facta est nubes et obumbrauit eos et timuerunt intrantibus il-

lis in nubem 9:35 et uox facta est de nube dicens hic est 9:35 And there came a voice out of the cloud, saying, This

is my beloved Son: hear him.

Filius meus electus ipsum audite 9:36 et dum fieret uox in- 9:36 And when the voice was past, Jesus was found alone.

And they kept it close, and told no man in those days any of those things which they had seen.

uentus est Iesus solus et ipsi tacuerunt et nemini dixerunt in illis diebus quicquam ex his quae uiderant 9:37 factum 9:37 And it came to pass, that on the next day, when they

were come down from the hill, much people met him.

est autem in sequenti die descendentibus illis de monte occurrit illi turba multa 9:38 et ecce uir de turba exclamauit 9:38 And, behold, a man of the company cried out, saying,

Master, I beseech thee, look upon my son: for he is mine only child.

dicens magister obsecro te respice in filium meum quia unicus est mihi 9:39 et ecce spiritus adprehendit illum et su- 9:39 And, lo, a spirit taketh him, and he suddenly crieth

out; and it teareth him that he foameth again, and bruising him hardly departeth from him.

bito clamat et elidit et dissipat eum cum spuma et uix discedit dilanians eum 9:40 et rogaui discipulos tuos ut ei- 9:40 And I besought thy disciples to cast him out; and they

could not.

cerent illum et non potuerunt 9:41 respondens autem Iesus 9:41 And Jesus answering said, O faithless and perverse

generation, how long shall I be with you, and suffer you? Bring thy son hither.

dixit o generatio infidelis et peruersa usquequo ero apud uos et patiar uos adduc huc filium tuum 9:42 et cum accede- 9:42 And as he was yet a coming, the devil threw him

down, and tare him. And Jesus rebuked the unclean spirit, and healed the child, and delivered him again to his father.

ret elisit illum daemonium et dissipauit 9:43 et increpauit

9:43 And they were all amazed at the mighty power of God. But while they wondered every one at all things which Jesus did, he said unto his disciples,

Iesus spiritum inmundum et sanauit puerum et reddidit illum patri eius 9:44 stupebant autem omnes in magnitudine 9:44 Let these sayings sink down into your ears: for the

Son of man shall be delivered into the hands of men.

Dei omnibusque mirantibus in omnibus quae faciebat dixit ad discipulos suos ponite uos in cordibus uestris sermones istos Filius enim hominis futurum est ut tradatur in manus hominum 9:45 at illi ignorabant uerbum istud et erat uela- 9:45 But they understood not this saying, and it was hid

from them, that they perceived it not: and they feared to ask him of that saying.

tum ante eos ut non sentirent illud et timebant interrogare eum de hoc uerbo 9:46 intrauit autem cogitatio in eos quis 9:46 Then there arose a reasoning among them, which of

them should be greatest.

eorum maior esset 9:47 at Iesus uidens cogitationes cordis 9:47 And Jesus, perceiving the thought of their heart, took

a child, and set him by him,

illorum adprehendens puerum statuit eum secus se 9:48 et 9:48 And said unto them, Whosoever shall receive this

child in my name receiveth me: and whosoever shall receive me receiveth him that sent me: for he that is least

ait illis quicumque susceperit puerum istum in nomine meo among you all, the same shall be great. me recipit et quicumque me recipit recipit eum qui me misit nam qui minor est inter omnes uos hic maior est 9:49 res- 9:49 And John answered and said, Master, we saw one

casting out devils in thy name; and we forbad him, because he followeth not with us.

Biblia Sacra Vulgata

pondens autem Iohannes dixit praeceptor uidimus quendam in nomine tuo eicientem daemonia et prohibuimus eum quia

9:50 And Jesus said unto him, Forbid him not: for he that non sequitur nobiscum 9:50 et ait ad illum Iesus nolite pro-

is not against us is for us.

9:51 And it came to pass, when the time was come that he hibere qui enim non est aduersum uos pro uobis est 9:51

should be received up, he stedfastly set his face to go to Jerusalem,

factum est autem dum conplerentur dies adsumptionis eius

9:52 And sent messengers before his face: and they went, et ipse faciem suam firmauit ut iret Hierusalem 9:52 et mi-

and entered into a village of the Samaritans, to make ready for him.

sit nuntios ante conspectum suum et euntes intrauerunt in ci-

9:53 And they did not receive him, because his face was uitatem Samaritanorum ut pararent illi 9:53 et non recepe-

as though he would go to Jerusalem.

9:54 And when his disciples James and John saw this, they runt eum quia facies eius erat euntis Hierusalem 9:54 cum

said, Lord, wilt thou that we command fire to come down from heaven, and consume them, even as Elias did?

uidissent autem discipuli eius Iacobus et Iohannes dixerunt Domine uis dicimus ut ignis descendat de caelo et consumat

9:55 But he turned, and rebuked them, and said, Ye know illos 9:55 et conuersus increpauit illos 9:56 et abierunt in

not what manner of spirit ye are of. 9:56 For the Son of man is not come to destroy men’s lives, but to save them. And they went to another village.

aliud castellum 9:57 factum est autem ambulantibus illis in

9:57 And it came to pass, that, as they went in the way, a certain man said unto him, Lord, I will follow thee whithersoever thou goest.

uia dixit quidam ad illum sequar te quocumque ieris 9:58

9:58 And Jesus said unto him, Foxes have holes, and birds of the air have nests; but the Son of man hath not where to lay his head.

et ait illi Iesus uulpes foueas habent et uolucres caeli nidos Filius autem hominis non habet ubi caput reclinet 9:59 ait

9:59 And he said unto another, Follow me. But he said, Lord, suffer me first to go and bury my father.

autem ad alterum sequere me ille autem dixit Domine per-

9:60 Jesus said unto him, Let the dead bury their dead: but mitte mihi primum ire sepelire patrem meum 9:60 dixitque go thou and preach the kingdom of God.

ei Iesus sine ut mortui sepeliant mortuos suos tu autem uade

9:61 And another also said, Lord, I will follow thee; but adnuntia regnum Dei 9:61 et ait alter sequar te Domine sed

let me first go bid them farewell, which are at home at my house.

primum permitte mihi renuntiare his qui domi sunt 9:62 ait

9:62 And Jesus said unto him, No man, having put his hand to the plough, and looking back, is fit for the kingdom of God.

ad illum Iesus nemo mittens manum suam in aratrum et aspiciens retro aptus est regno Dei

10:1 After these things the LORD appointed other seventy 10:1 post haec autem designauit Dominus et alios septuaalso, and sent them two and two before his face into every city and place, whither he himself would come.

ginta duos et misit illos binos ante faciem suam in omnem

10:2 Therefore said he unto them, The harvest truly is ciuitatem et locum quo erat ipse uenturus 10:2 et dicebat il-

great, but the labourers are few: pray ye therefore the Lord of the harvest, that he would send forth labourers into his

harvest. lis messis quidem multa operarii autem pauci rogate ergo 10:3 Go your ways: behold, I send you forth as lambs Dominum messis ut mittat operarios in messem 10:3 ite

among wolves. 10:4 Carry neither purse, nor scrip, nor shoes: and salute ecce ego mitto uos sicut agnos inter lupos 10:4 nolite porno man by the way.

tare sacculum neque peram neque calciamenta et neminem

10:5 And into whatsoever house ye enter, first say, Peace per uiam salutaueritis 10:5 in quamcumque domum intrabe to this house. 10:6 And if the son of peace be there, your peace shall rest ueritis primum dicite pax huic domui 10:6 et si ibi fuerit upon it: if not, it shall turn to you again.

filius pacis requiescet super illam pax uestra sin autem ad

Euangelium Secundum Lucam

uos reuertetur 10:7 in eadem autem domo manete edentes 10:7 And in the same house remain, eating and drinking

such things as they give: for the labourer is worthy of his hire. Go not from house to house.

et bibentes quae apud illos sunt dignus enim est operarius

mercede sua nolite transire de domo in domum 10:8 et in 10:8 And into whatsoever city ye enter, and they receive

you, eat such things as are set before you:

quamcumque ciuitatem intraueritis et susceperint uos man-

ducate quae adponuntur uobis 10:9 et curate infirmos qui 10:9 And heal the sick that are therein, and say unto them,

The kingdom of God is come nigh unto you.

in illa sunt et dicite illis adpropinquauit in uos regnum Dei 10:10 in quamcumque ciuitatem intraueritis et non receperint 10:10 But into whatsoever city ye enter, and they receive

you not, go your ways out into the streets of the same, and say,

uos exeuntes in plateas eius dicite 10:11 etiam puluerem

10:11 Even the very dust of your city, which cleaveth on us, we do wipe off against you: notwithstanding be ye sure of this, that the kingdom of God is come nigh unto you.

qui adhesit nobis de ciuitate uestra extergimus in uos tamen hoc scitote quia adpropinquauit regnum Dei 10:12 dico uo- 10:12 But I say unto you, that it shall be more tolerable in

that day for Sodom, than for that city.

bis quia Sodomis in die illa remissius erit quam illi ciuitati 10:13 uae tibi Corazain uae tibi Bethsaida quia si in Tyro et 10:13 Woe unto thee, Chorazin! woe unto thee, Bethsaida!

for if the mighty works had been done in Tyre and Sidon, which have been done in you, they had a great while ago

Sidone factae fuissent uirtutes quae in uobis factae sunt olim repented, sitting in sackcloth and ashes. in cilicio et cinere sedentes paeniterent 10:14 uerumtamen 10:14 But it shall be more tolerable for Tyre and Sidon at

the judgment, than for you.

Tyro et Sidoni remissius erit in iudicio quam uobis 10:15 10:15 And thou, Capernaum, which art exalted to heaven,

shalt be thrust down to hell.

et tu Capharnaum usque in caelum exaltata usque ad infernum demergeris 10:16 qui uos audit me audit et qui uos 10:16 He that heareth you heareth me; and he that de-

spiseth you despiseth me; and he that despiseth me despiseth him that sent me.

spernit me spernit qui autem me spernit spernit eum qui me misit 10:17 reuersi sunt autem septuaginta duo cum gaudio 10:17 And the seventy returned again with joy, saying,

Lord, even the devils are subject unto us through thy name.

dicentes Domine etiam daemonia subiciuntur nobis in nomine tuo 10:18 et ait illis uidebam Satanan sicut fulgur de 10:18 And he said unto them, I beheld Satan as lightning

fall from heaven.

caelo cadentem 10:19 ecce dedi uobis potestatem calcandi 10:19 Behold, I give unto you power to tread on serpents

and scorpions, and over all the power of the enemy: and nothing shall by any means hurt you.

supra serpentes et scorpiones et supra omnem uirtutem inimici et nihil uobis nocebit 10:20 uerumtamen in hoc nolite 10:20 Notwithstanding in this rejoice not, that the spir-

its are subject unto you; but rather rejoice, because your names are written in heaven.

gaudere quia spiritus uobis subiciuntur gaudete autem quod nomina uestra scripta sunt in caelis 10:21 in ipsa hora exul- 10:21 In that hour Jesus rejoiced in spirit, and said, I thank

thee, O Father, Lord of heaven and earth, that thou hast hid these things from the wise and prudent, and hast revealed them unto babes: even so, Father; for so it seemed good in thy sight.

tauit Spiritu Sancto et dixit confiteor tibi Pater Domine caeli et terrae quod abscondisti haec a sapientibus et prudentibus et reuelasti ea paruulis etiam Pater quia sic placuit ante te 10:22 omnia mihi tradita sunt a Patre meo et nemo scit qui 10:22 All things are delivered to me of my Father: and no

man knoweth who the Son is, but the Father; and who the Father is, but the Son, and he to whom the Son will reveal

sit Filius nisi Pater et qui sit Pater nisi Filius et cui uoluerit him. Filius reuelare 10:23 et conuersus ad discipulos suos dixit 10:23 And he turned him unto his disciples, and said pri-

vately, Blessed are the eyes which see the things that ye see:

beati oculi qui uident quae uidetis 10:24 dico enim uobis

10:24 For I tell you, that many prophets and kings have desired to see those things which ye see, and have not seen them; and to hear those things which ye hear, and have not heard them.

Biblia Sacra Vulgata

quod multi prophetae et reges uoluerunt uidere quae uos uidetis et non uiderunt et audire quae auditis et non audierunt

10:25 And, behold, a certain lawyer stood up, and tempted 10:25 et ecce quidam legis peritus surrexit temptans illum

him, saying, Master, what shall I do to inherit eternal life?

et dicens magister quid faciendo uitam aeternam possidebo

10:26 He said unto him, What is written in the law? how 10:26 at ille dixit ad eum in lege quid scriptum est quomodo

readest thou? 10:27 And he answering said, Thou shalt love the Lord thy legis 10:27 ille respondens dixit diliges Dominum Deum God with all thy heart, and with all thy soul, and with all thy strength, and with all thy mind; and thy neighbour as

thyself. tuum ex toto corde tuo et ex tota anima tua et ex omnibus

uiribus tuis et ex omni mente tua et proximum tuum sic-

10:28 And he said unto him, Thou hast answered right: ut te ipsum 10:28 dixitque illi recte respondisti hoc fac et

this do, and thou shalt live. 10:29 But he, willing to justify himself, said unto Jesus, uiues 10:29 ille autem uolens iustificare se ipsum dixit ad And who is my neighbour? 10:30 And Jesus answering said, A certain man went down Iesum et quis est meus proximus 10:30 suscipiens autem from Jerusalem to Jericho, and fell among thieves, which stripped him of his raiment, and wounded him, and de-

parted, leaving him half dead. Iesus dixit homo quidam descendebat ab Hierusalem in Hie-

richo et incidit in latrones qui etiam despoliauerunt eum et

10:31 And by chance there came down a certain priest that plagis inpositis abierunt semiuiuo relicto 10:31 accidit au-

way: and when he saw him, he passed by on the other side.

tem ut sacerdos quidam descenderet eadem uia et uiso illo

10:32 And likewise a Levite, when he was at the place, praeteriuit 10:32 similiter et Leuita cum esset secus locum

came and looked on him, and passed by on the other side.

10:33 But a certain Samaritan, as he journeyed, came et uideret eum pertransiit 10:33 Samaritanus autem quidam

where he was: and when he saw him, he had compassion on him,

iter faciens uenit secus eum et uidens eum misericordia mo-

10:34 And went to him, and bound up his wounds, pouring tus est 10:34 et adpropians alligauit uulnera eius infundens

in oil and wine, and set him on his own beast, and brought him to an inn, and took care of him.

oleum et uinum et inponens illum in iumentum suum duxit in

10:35 And on the morrow when he departed, he took out stabulum et curam eius egit 10:35 et altera die protulit duos

two pence, and gave them to the host, and said unto him, Take care of him; and whatsoever thou spendest more,

when I come again, I will repay thee. denarios et dedit stabulario et ait curam illius habe et quod-10:36 Which now of these three, thinkest thou, was neigh- cumque supererogaueris ego cum rediiro reddam tibi 10:36

bour unto him that fell among the thieves?

quis horum trium uidetur tibi proximus fuisse illi qui incidit

10:37 And he said, He that shewed mercy on him. Then in latrones 10:37 at ille dixit qui fecit misericordiam in il-

said Jesus unto him, Go, and do thou likewise.

10:38 Now it came to pass, as they went, that he entered lum et ait illi Iesus uade et tu fac similiter 10:38 factum est

into a certain village: and a certain woman named Martha received him into her house.

autem dum irent et ipse intrauit in quoddam castellum et mulier quaedam Martha nomine excepit illum in domum suam

10:39 And she had a sister called Mary, which also sat at 10:39 et huic erat soror nomine Maria quae etiam sedens se-

Jesus’ feet, and heard his word.

10:40 But Martha was cumbered about much serving, and cus pedes Domini audiebat uerbum illius 10:40 Martha au-

came to him, and said, Lord, dost thou not care that my sister hath left me to serve alone? bid her therefore that

she help me. tem satagebat circa frequens ministerium quae stetit et ait

Domine non est tibi curae quod soror mea reliquit me solam

10:41 And Jesus answered and said unto her, Martha, ministrare dic ergo illi ut me adiuuet 10:41 et respondens

Martha, thou art careful and troubled about many things:

Euangelium Secundum Lucam

dixit illi Dominus Martha Martha sollicita es et turbaris erga plurima 10:42 porro unum est necessarium Maria optimam 10:42 But one thing is needful: and Mary hath chosen that

good part, which shall not be taken away from her.

partem elegit quae non auferetur ab ea

11:1 et factum est cum esset in loco quodam orans ut ces- 11:1 And it came to pass, that, as he was praying in a

certain place, when he ceased, one of his disciples said unto him, Lord, teach us to pray, as John also taught his

sauit dixit unus ex discipulis eius ad eum Domine doce nos disciples. orare sicut et Iohannes docuit discipulos suos 11:2 et ait il- 11:2 And he said unto them, When ye pray, say, Our Father

which art in heaven, Hallowed be thy name. Thy kingdom come. Thy will be done, as in heaven, so in earth.

lis cum oratis dicite Pater sanctificetur nomen tuum adueniat regnum tuum 11:3 panem nostrum cotidianum da nobis co- 11:3 Give us day by day our daily bread. tidie 11:4 et dimitte nobis peccata nostra siquidem et ipsi 11:4 And forgive us our sins; for we also forgive every one

that is indebted to us. And lead us not into temptation; but deliver us from evil.

dimittimus omni debenti nobis et ne nos inducas in temptationem 11:5 et ait ad illos quis uestrum habebit amicum et 11:5 And he said unto them, Which of you shall have a

friend, and shall go unto him at midnight, and say unto him, Friend, lend me three loaves;

ibit ad illum media nocte et dicit illi amice commoda mihi tres panes 11:6 quoniam amicus meus uenit de uia ad me et 11:6 For a friend of mine in his journey is come to me, and

I have nothing to set before him?

non habeo quod ponam ante illum 11:7 et ille de intus re- 11:7 And he from within shall answer and say, Trouble me

not: the door is now shut, and my children are with me in bed; I cannot rise and give thee.

spondens dicat noli mihi molestus esse iam ostium clausum est et pueri mei mecum sunt in cubili non possum surgere et dare tibi 11:8 dico uobis et si non dabit illi surgens eo quod 11:8 I say unto you, Though he will not rise and give him,

because he is his friend, yet because of his importunity he will rise and give him as many as he needeth.

amicus eius sit propter inprobitatem tamen eius surget et dabit illi quotquot habet necessarios 11:9 et ego uobis dico 11:9 And I say unto you, Ask, and it shall be given you;

seek, and ye shall find; knock, and it shall be opened unto you.

petite et dabitur uobis quaerite et inuenietis pulsate et aperietur uobis 11:10 omnis enim qui petit accipit et qui quae- 11:10 For every one that asketh receiveth; and he that

seeketh findeth; and to him that knocketh it shall be opened.

rit inuenit et pulsanti aperietur 11:11 quis autem ex uobis

11:11 If a son shall ask bread of any of you that is a father, will he give him a stone? or if he ask a fish, will he for a fish give him a serpent?

patrem petet panem numquid lapidem dabit illi aut piscem numquid pro pisce serpentem dabit illi 11:12 aut si petierit 11:12 Or if he shall ask an egg, will he offer him a scor-

pion?

ouum numquid porriget illi scorpionem 11:13 si ergo uos 11:13 If ye then, being evil, know how to give good gifts

unto your children: how much more shall your heavenly Father give the Holy Spirit to them that ask him?

cum sitis mali nostis bona data dare filiis uestris quanto magis Pater uester de caelo dabit spiritum bonum petentibus se 11:14 et erat eiciens daemonium et illud erat mutum et cum 11:14 And he was casting out a devil, and it was dumb.

And it came to pass, when the devil was gone out, the dumb spake; and the people wondered.

eiecisset daemonium locutus est mutus et admiratae sunt turbae 11:15 quidam autem ex eis dixerunt in Beelzebub prin- 11:15 But some of them said, He casteth out devils

through Beelzebub the chief of the devils.

cipe daemoniorum eicit daemonia 11:16 et alii temptantes 11:16 And others, tempting him, sought of him a sign from

heaven.

signum de caelo quaerebant ab eo 11:17 ipse autem ut uidit 11:17 But he, knowing their thoughts, said unto them, Ev-

ery kingdom divided against itself is brought to desolation; and a house divided against a house falleth.

cogitationes eorum dixit eis omne regnum in se ipsum diui-

Biblia Sacra Vulgata

11:18 If Satan also be divided against himself, how shall sum desolatur et domus supra domum cadet 11:18 si autem his kingdom stand? because ye say that I cast out devils through Beelzebub.

et Satanas in se ipsum diuisus est quomodo stabit regnum

11:19 And if I by Beelzebub cast out devils, by whom ipsius quia dicitis in Beelzebub eicere me daemonia 11:19

do your sons cast them out? therefore shall they be your judges.

si autem ego in Beelzebub eicio daemonia filii uestri in quo

11:20 But if I with the finger of God cast out devils, no eiciunt ideo ipsi iudices uestri erunt 11:20 porro si in digito

doubt the kingdom of God is come upon you.

Dei eicio daemonia profecto praeuenit in uos regnum Dei

11:21 When a strong man armed keepeth his palace, his 11:21 cum fortis armatus custodit atrium suum in pace sunt

goods are in peace:

11:22 But when a stronger than he shall come upon him, ea quae possidet 11:22 si autem fortior illo superueniens ui-

and overcome him, he taketh from him all his armour wherein he trusted, and divideth his spoils.

cerit eum uniuersa arma eius aufert in quibus confidebat et

11:23 He that is not with me is against me: and he that spolia eius distribuit 11:23 qui non est mecum aduersum me

gathereth not with me scattereth.

11:24 When the unclean spirit is gone out of a man, est et qui non colligit mecum dispergit 11:24 cum inmun-

he walketh through dry places, seeking rest; and finding none, he saith, I will return unto my house whence I came

out. dus spiritus exierit de homine perambulat per loca inaquosa

quaerens requiem et non inueniens dicit reuertar in domum

11:25 And when he cometh, he findeth it swept and gar- meam unde exiui 11:25 et cum uenerit inuenit scopis mun-

nished.

11:26 Then goeth he, and taketh to him seven other spirits datam 11:26 et tunc uadit et adsumit septem alios spiritus

more wicked than himself; and they enter in, and dwell there: and the last state of that man is worse than the first.

nequiores se et ingressi habitant ibi et sunt nouissima ho-

11:27 And it came to pass, as he spake these things, a minis illius peiora prioribus 11:27 factum est autem cum

certain woman of the company lifted up her voice, and said unto him, Blessed is the womb that bare thee, and the

paps which thou hast sucked. haec diceret extollens uocem quaedam mulier de turba dixit 11:28 But he said, Yea rather, blessed are they that hear illi beatus uenter qui te portauit et ubera quae suxisti 11:28

the word of God, and keep it.

at ille dixit quippini beati qui audiunt uerbum Dei et custo-

11:29 And when the people were gathered thick together, diunt 11:29 turbis autem concurrentibus coepit dicere ge-

he began to say, This is an evil generation: they seek a sign; and there shall no sign be given it, but the sign of

Jonas the prophet. neratio haec generatio nequam est signum quaerit et signum 11:30 For as Jonas was a sign unto the Ninevites, so shall non dabitur illi nisi signum Ionae 11:30 nam sicut Ionas fuit

also the Son of man be to this generation.

signum Nineuitis ita erit et Filius hominis generationi isti

11:31 The queen of the south shall rise up in the judgment 11:31 regina austri surget in iudicio cum uiris generationis

with the men of this generation, and condemn them: for she came from the utmost parts of the earth to hear the wisdom of Solomon; and, behold, a greater than Solomon is here.

huius et condemnabit illos quia uenit a finibus terrae audire

11:32 The men of Nineve shall rise up in the judgment sapientiam Salomonis et ecce plus Salomone hic 11:32 uiri

with this generation, and shall condemn it: for they repented at the preaching of Jonas; and, behold, a greater

than Jonas is here. nineuitae surgent in iudicio cum generatione hac et condem-

nabunt illam quia paenitentiam egerunt ad praedicationem

11:33 No man, when he hath lighted a candle, putteth it in Ionae et ecce plus Iona hic 11:33 nemo lucernam accendit a secret place, neither under a bushel, but on a candlestick, that they which come in may see the light.

et in abscondito ponit neque sub modio sed supra candela-

11:34 The light of the body is the eye: therefore when brum ut qui ingrediuntur lumen uideant 11:34 lucerna cor-

thine eye is single, thy whole body also is full of light; but when thine eye is evil, thy body also is full of darkness.

poris tui est oculus tuus si oculus tuus fuerit simplex totum

Euangelium Secundum Lucam

corpus tuum lucidum erit si autem nequam fuerit etiam cor-

pus tuum tenebrosum erit 11:35 uide ergo ne lumen quod 11:35 Take heed therefore that the light which is in thee be

not darkness.

in te est tenebrae sint 11:36 si ergo corpus tuum totum luci- 11:36 If thy whole body therefore be full of light, having

no part dark, the whole shall be full of light, as when the bright shining of a candle doth give thee light.

dum fuerit non habens aliquam partem tenebrarum erit lucidum totum et sicut lucerna fulgoris inluminabit te 11:37 et 11:37 And as he spake, a certain Pharisee besought him to

dine with him: and he went in, and sat down to meat.

cum loqueretur rogauit illum quidam Pharisaeus ut pranderet apud se et ingressus recubuit 11:38 Pharisaeus autem coe- 11:38 And when the Pharisee saw it, he marvelled that he

had not first washed before dinner.

pit intra se reputans dicere quare non baptizatus esset ante prandium 11:39 et ait Dominus ad illum nunc uos Pharisaei 11:39 And the Lord said unto him, Now do ye Pharisees

make clean the outside of the cup and the platter; but your inward part is full of ravening and wickedness.

quod de foris est calicis et catini mundatis quod autem intus est uestrum plenum est rapina et iniquitate 11:40 stulti 11:40 Ye fools, did not he that made that which is without

make that which is within also?

nonne qui fecit quod de foris est etiam id quod de intus est fecit 11:41 uerumtamen quod superest date elemosynam et 11:41 But rather give alms of such things as ye have; and,

behold, all things are clean unto you.

ecce omnia munda sunt uobis 11:42 sed uae uobis Phari- 11:42 But woe unto you, Pharisees! for ye tithe mint and rue and all manner of herbs, and pass over judgment and the love of God: these ought ye to have done, and not to

saeis quia decimatis mentam et rutam et omne holus et prae- leave the other undone. teritis iudicium et caritatem Dei haec autem oportuit facere et illa non omittere 11:43 uae uobis Pharisaeis quia diligitis 11:43 Woe unto you, Pharisees! for ye love the uppermost

seats in the synagogues, and greetings in the markets.

primas cathedras in synagogis et salutationes in foro 11:44 11:44 Woe unto you, scribes and Pharisees, hypocrites!

for ye are as graves which appear not, and the men that walk over them are not aware of them.

uae uobis quia estis ut monumenta quae non parent et homines ambulantes supra nesciunt 11:45 respondens autem qui- 11:45 Then answered one of the lawyers, and said unto

him, Master, thus saying thou reproachest us also.

dam ex legis peritis ait illi magister haec dicens etiam nobis contumeliam facis 11:46 at ille ait et uobis legis peritis uae 11:46 And he said, Woe unto you also, ye lawyers! for

ye lade men with burdens grievous to be borne, and ye yourselves touch not the burdens with one of your fingers.

quia oneratis homines oneribus quae portari non possunt et ipsi uno digito uestro non tangitis sarcinas 11:47 uae uobis 11:47 Woe unto you! for ye build the sepulchres of the

prophets, and your fathers killed them.

quia aedificatis monumenta prophetarum patres autem uestri occiderunt illos 11:48 profecto testificamini quod consenti- 11:48 Truly ye bear witness that ye allow the deeds of your

fathers: for they indeed killed them, and ye build their sepulchres.

tis operibus patrum uestrorum quoniam quidem ipsi eos occiderunt uos autem aedificatis eorum sepulchra 11:49 prop- 11:49 Therefore also said the wisdom of God, I will send

them prophets and apostles, and some of them they shall slay and persecute:

terea et sapientia Dei dixit mittam ad illos prophetas et apostolos et ex illis occident et persequentur 11:50 ut inquiratur 11:50 That the blood of all the prophets, which was shed

from the foundation of the world, may be required of this generation;

sanguis omnium prophetarum qui effusus est a constitutione mundi a generatione ista 11:51 a sanguine Abel usque ad 11:51 From the blood of Abel unto the blood of Zacharias

which perished between the altar and the temple: verily I say unto you, It shall be required of this generation.

sanguinem Zacchariae qui periit inter altare et aedem ita dico

Biblia Sacra Vulgata

11:52 Woe unto you, lawyers! for ye have taken away the uobis requiretur ab hac generatione 11:52 uae uobis legis

key of knowledge: ye entered not in yourselves, and them that were entering in ye hindered.

peritis quia tulistis clauem scientiae ipsi non introistis et eos

11:53 And as he said these things unto them, the scribes qui introibant prohibuistis 11:53 cum haec ad illos diceret

and the Pharisees began to urge him vehemently, and to provoke him to speak of many things:

coeperunt Pharisaei et legis periti grauiter insistere et os eius

11:54 Laying wait for him, and seeking to catch something opprimere de multis 11:54 insidiantes et quaerentes capere

out of his mouth, that they might accuse him.

aliquid ex ore eius ut accusarent eum

12:1 In the mean time, when there were gathered together 12:1 multis autem turbis circumstantibus ita ut se inuicem

an innumerable multitude of people, insomuch that they trode one upon another, he began to say unto his disciples first of all, Beware ye of the leaven of the Pharisees, which is hypocrisy.

conculcarent coepit dicere ad discipulos suos adtendite a fer-

12:2 For there is nothing covered, that shall not be re- mento Pharisaeorum quae est hypocrisis 12:2 nihil autem

vealed; neither hid, that shall not be known.

opertum est quod non reueletur neque absconditum quod

12:3 Therefore whatsoever ye have spoken in darkness non sciatur 12:3 quoniam quae in tenebris dixistis in lumine

shall be heard in the light; and that which ye have spoken in the ear in closets shall be proclaimed upon the house-

tops. dicentur et quod in aurem locuti estis in cubiculis praedica-12:4 And I say unto you my friends, Be not afraid of them bitur in tectis 12:4 dico autem uobis amicis meis ne terre-

that kill the body, and after that have no more that they can do.

amini ab his qui occidunt corpus et post haec non habent

12:5 But I will forewarn you whom ye shall fear: Fear him, amplius quod faciant 12:5 ostendam autem uobis quem tiwhich after he hath killed hath power to cast into hell; yea, I say unto you, Fear him.

meatis timete eum qui postquam occiderit habet potestatem

12:6 Are not five sparrows sold for two farthings, and not mittere in gehennam ita dico uobis hunc timete 12:6 nonne

one of them is forgotten before God?

quinque passeres ueneunt dipundio et unus ex illis non est in

12:7 But even the very hairs of your head are all num- obliuione coram Deo 12:7 sed et capilli capitis uestri om-

bered. Fear not therefore: ye are of more value than many sparrows.

nes numerati sunt nolite ergo timere multis passeribus plu-

12:8 Also I say unto you, Whosoever shall confess me ris estis 12:8 dico autem uobis omnis quicumque confessus

before men, him shall the Son of man also confess before the angels of God:

fuerit in me coram hominibus et Filius hominis confitebitur

12:9 But he that denieth me before men shall be denied in illo coram angelis Dei 12:9 qui autem negauerit me co-

before the angels of God.

12:10 And whosoever shall speak a word against the Son ram hominibus denegabitur coram angelis Dei 12:10 et om-

of man, it shall be forgiven him: but unto him that blasphemeth against the Holy Ghost it shall not be forgiven.

nis qui dicit uerbum in Filium hominis remittetur illi ei autem qui in Spiritum Sanctum blasphemauerit non remittetur

12:11 And when they bring you unto the synagogues, and 12:11 cum autem inducent uos in synagogas et ad magistratus

unto magistrates, and powers, take ye no thought how or what thing ye shall answer, or what ye shall say:

et potestates nolite solliciti esse qualiter aut quid respondea-

12:12 For the Holy Ghost shall teach you in the same hour tis aut quid dicatis 12:12 Spiritus enim Sanctus docebit uos

what ye ought to say.

12:13 And one of the company said unto him, Master, in ipsa hora quae oporteat dicere 12:13 ait autem quidam ei

speak to my brother, that he divide the inheritance with me.

de turba magister dic fratri meo ut diuidat mecum heredita-

12:14 And he said unto him, Man, who made me a judge tem 12:14 at ille dixit ei homo quis me constituit iudicem

or a divider over you?

12:15 And he said unto them, Take heed, and beware of aut diuisorem super uos 12:15 dixitque ad illos uidete et ca-

covetousness: for a man’s life consisteth not in the abundance of the things which he possesseth.

Euangelium Secundum Lucam

uete ab omni auaritia quia non in abundantia cuiusquam uita eius est ex his quae possidet 12:16 dixit autem similitudi- 12:16 And he spake a parable unto them, saying, The

ground of a certain rich man brought forth plentifully:

nem ad illos dicens hominis cuiusdam diuitis uberes fructus ager adtulit 12:17 et cogitabat intra se dicens quid faciam 12:17 And he thought within himself, saying, What shall

I do, because I have no room where to bestow my fruits?

quod non habeo quo congregem fructus meos 12:18 et dixit 12:18 And he said, This will I do: I will pull down my

barns, and build greater; and there will I bestow all my fruits and my goods.

hoc faciam destruam horrea mea et maiora faciam et illuc congregabo omnia quae nata sunt mihi et bona mea 12:19 12:19 And I will say to my soul, Soul, thou hast much

goods laid up for many years; take thine ease, eat, drink, and be merry.

et dicam animae meae anima habes multa bona posita in annos plurimos requiesce comede bibe epulare 12:20 dixit au- 12:20 But God said unto him, Thou fool, this night thy

soul shall be required of thee: then whose shall those things be, which thou hast provided?

tem illi Deus stulte hac nocte animam tuam repetunt a te quae autem parasti cuius erunt 12:21 sic est qui sibi thesau- 12:21 So is he that layeth up treasure for himself, and is

not rich toward God.

rizat et non est in Deum diues 12:22 dixitque ad discipulos 12:22 And he said unto his disciples, Therefore I say unto

you, Take no thought for your life, what ye shall eat; neither for the body, what ye shall put on.

suos ideo dico uobis nolite solliciti esse animae quid manducetis neque corpori quid uestiamini 12:23 anima plus est 12:23 The life is more than meat, and the body is more

than raiment.

quam esca et corpus quam uestimentum 12:24 considerate 12:24 Consider the ravens: for they neither sow nor reap; which neither have storehouse nor barn; and God feedeth them: how much more are ye better than the fowls?

coruos quia non seminant neque metunt quibus non est cellarium neque horreum et Deus pascit illos quanto magis uos pluris estis illis 12:25 quis autem uestrum cogitando potest 12:25 And which of you with taking thought can add to

his stature one cubit?

adicere ad staturam suam cubitum unum 12:26 si ergo ne- 12:26 If ye then be not able to do that thing which is least,

why take ye thought for the rest?

que quod minimum est potestis quid de ceteris solliciti estis 12:27 considerate lilia quomodo crescunt non laborant non 12:27 Consider the lilies how they grow: they toil not,

they spin not; and yet I say unto you, that Solomon in all his glory was not arrayed like one of these.

nent dico autem uobis nec Salomon in omni gloria sua uestiebatur sicut unum ex istis 12:28 si autem faenum quod 12:28 If then God so clothe the grass, which is to day in

the field, and to morrow is cast into the oven; how much more will he clothe you, O ye of little faith?

hodie in agro est et cras in clibanum mittitur Deus sic uestit quanto magis uos pusillae fidei 12:29 et uos nolite quae- 12:29 And seek not ye what ye shall eat, or what ye shall

drink, neither be ye of doubtful mind.

rere quid manducetis aut quid bibatis et nolite in sublime

tolli 12:30 haec enim omnia gentes mundi quaerunt Pater 12:30 For all these things do the nations of the world seek after: and your Father knoweth that ye have need of these things.

autem uester scit quoniam his indigetis 12:31 uerumtamen

12:31 But rather seek ye the kingdom of God; and all these things shall be added unto you.

quaerite regnum Dei et haec omnia adicientur uobis 12:32 _{12:32 Fear not, little flock; for it is your Father’s good}

pleasure to give you the kingdom.

nolite timere pusillus grex quia conplacuit Patri uestro dare uobis regnum 12:33 uendite quae possidetis et date elemo- 12:33 Sell that ye have, and give alms; provide yourselves

bags which wax not old, a treasure in the heavens that faileth not, where no thief approacheth, neither moth cor-

synam facite uobis sacculos qui non ueterescunt thesaurum rupteth. non deficientem in caelis quo fur non adpropiat neque ti-

Biblia Sacra Vulgata

12:34 For where your treasure is, there will your heart be nea corrumpit 12:34 ubi enim thesaurus uester est ibi et cor

also.

12:35 Let your loins be girded about, and your lights burn- uestrum erit 12:35 sint lumbi uestri praecincti et lucernae

ing;

12:36 And ye yourselves like unto men that wait for their ardentes 12:36 et uos similes hominibus expectantibus do-

lord, when he will return from the wedding; that when he cometh and knocketh, they may open unto him immedi-

ately. minum suum quando reuertatur a nuptiis ut cum uenerit et 12:37 Blessed are those servants, whom the lord when he pulsauerit confestim aperiant ei 12:37 beati serui illi quos

cometh shall find watching: verily I say unto you, that he shall gird himself, and make them to sit down to meat, and

will come forth and serve them. cum uenerit dominus inuenerit uigilantes amen dico uobis

quod praecinget se et faciet illos discumbere et transiens

12:38 And if he shall come in the second watch, or come ministrabit illis 12:38 et si uenerit in secunda uigilia et si

in the third watch, and find them so, blessed are those servants.

in tertia uigilia uenerit et ita inuenerit beati sunt serui illi

12:39 And this know, that if the goodman of the house had 12:39 hoc autem scitote quia si sciret pater familias qua hora

known what hour the thief would come, he would have watched, and not have suffered his house to be broken

through. fur ueniret uigilaret utique et non sineret perfodiri domum 12:40 Be ye therefore ready also: for the Son of man suam 12:40 et uos estote parati quia qua hora non putatis

cometh at an hour when ye think not. 12:41 Then Peter said unto him, Lord, speakest thou this Filius hominis uenit 12:41 ait autem ei Petrus Domine ad parable unto us, or even to all?

12:42 And the Lord said, Who then is that faithful and nos dicis hanc parabolam an et ad omnes 12:42 dixit autem

wise steward, whom his lord shall make ruler over his household, to give them their portion of meat in due sea-

son? Dominus quis putas est fidelis dispensator et prudens quem

constituet dominus super familiam suam ut det illis in tem-

12:43 Blessed is that servant, whom his lord when he pore tritici mensuram 12:43 beatus ille seruus quem cum

cometh shall find so doing.

12:44 Of a truth I say unto you, that he will make him ruler uenerit dominus inuenerit ita facientem 12:44 uere dico uo-

over all that he hath.

12:45 But and if that servant say in his heart, My lord de- bis quia supra omnia quae possidet constituet illum 12:45

layeth his coming; and shall begin to beat the menservants and maidens, and to eat and drink, and to be drunken;

quod si dixerit seruus ille in corde suo moram facit dominus meus uenire et coeperit percutere pueros et ancillas et

12:46 The lord of that servant will come in a day when he edere et bibere et inebriari 12:46 ueniet dominus serui il-

looketh not for him, and at an hour when he is not aware, and will cut him in sunder, and will appoint him his por-

tion with the unbelievers. lius in die qua non sperat et hora qua nescit et diuidet eum 12:47 And that servant, which knew his lord’s will, and partemque eius cum infidelibus ponet 12:47 ille autem ser-

prepared not himself, neither did according to his will, shall be beaten with many stripes.

uus qui cognouit uoluntatem domini sui et non praeparauit et

12:48 But he that knew not, and did commit things wor- non fecit secundum uoluntatem eius uapulabit multas 12:48

thy of stripes, shall be beaten with few stripes. For unto whomsoever much is given, of him shall be much required: and to whom men have committed much, of him they will ask the more.

qui autem non cognouit et fecit digna plagis uapulabit paucis omni autem cui multum datum est multum quaeretur ab eo

12:49 I am come to send fire on the earth; and what will I, et cui commendauerunt multum plus petent ab eo 12:49 ig-

if it be already kindled?

12:50 But I have a baptism to be baptized with; and how nem ueni mittere in terram et quid uolo si accendatur 12:50

am I straitened till it be accomplished!

baptisma autem habeo baptizari et quomodo coartor usque

12:51 Suppose ye that I am come to give peace on earth? dum perficiatur 12:51 putatis quia pacem ueni dare in ter-

I tell you, Nay; but rather division: 12:52 For from henceforth there shall be five in one house ram non dico uobis sed separationem 12:52 erunt enim ex divided, three against two, and two against three.

Euangelium Secundum Lucam

hoc quinque in domo una diuisi tres in duo et duo in tres 12:53 diuidentur pater in filium et filius in patrem suum ma- 12:53 The father shall be divided against the son, and the

son against the father; the mother against the daughter, and the daughter against the mother; the mother in law against her daughter in law, and the daughter in law against her mother in law.

ter in filiam et filia in matrem socrus in nurum suam et nurus in socrum suam 12:54 dicebat autem et ad turbas cum uide- 12:54 And he said also to the people, When ye see a cloud

rise out of the west, straightway ye say, There cometh a shower; and so it is.

ritis nubem orientem ab occasu statim dicitis nimbus uenit et ita fit 12:55 et cum austrum flantem dicitis quia aestus erit 12:55 And when ye see the south wind blow, ye say, There

will be heat; and it cometh to pass.

et fit 12:56 hypocritae faciem terrae et caeli nostis probare 12:56 Ye hypocrites, ye can discern the face of the sky and

of the earth; but how is it that ye do not discern this time?

hoc autem tempus quomodo non probatis 12:57 quid autem 12:57 Yea, and why even of yourselves judge ye not what

is right?

et a uobis ipsis non iudicatis quod iustum est 12:58 cum au- 12:58 When thou goest with thine adversary to the magis-

trate, as thou art in the way, give diligence that thou mayest be delivered from him; lest he hale thee to the judge, and the judge deliver thee to the officer, and the officer cast thee into prison.

tem uadis cum aduersario tuo ad principem in uia da operam liberari ab illo ne forte trahat te apud iudicem et iudex tradat

te exactori et exactor mittat te in carcerem 12:59 dico tibi 12:59 I tell thee, thou shalt not depart thence, till thou hast

paid the very last mite.

non exies inde donec etiam nouissimum minutum reddas

13:1 aderant autem quidam ipso in tempore nuntiantes illi 13:1 There were present at that season some that told him

of the Galilaeans, whose blood Pilate had mingled with their sacrifices.

de Galilaeis quorum sanguinem Pilatus miscuit cum sacrificiis eorum 13:2 et respondens dixit illis putatis quod hii 13:2 And Jesus answering said unto them, Suppose ye that

these Galilaeans were sinners above all the Galilaeans, because they suffered such things?

Galilaei prae omnibus Galilaeis peccatores fuerunt quia talia passi sunt 13:3 non dico uobis sed nisi paenitentiam habue- 13:3 I tell you, Nay: but, except ye repent, ye shall all

likewise perish.

ritis omnes similiter peribitis 13:4 sicut illi decem et octo 13:4 Or those eighteen, upon whom the tower in Siloam

fell, and slew them, think ye that they were sinners above all men that dwelt in Jerusalem?

supra quos cecidit turris in Siloam et occidit eos putatis quia et ipsi debitores fuerunt praeter omnes homines habitantes in Hierusalem 13:5 non dico uobis sed si non paenitentiam 13:5 I tell you, Nay: but, except ye repent, ye shall all

likewise perish.

egeritis omnes similiter peribitis 13:6 dicebat autem hanc 13:6 He spake also this parable; A certain man had a fig

tree planted in his vineyard; and he came and sought fruit thereon, and found none.

similitudinem arborem fici habebat quidam plantatam in uinea sua et uenit quaerens fructum in illa et non inuenit 13:7 13:7 Then said he unto the dresser of his vineyard, Behold,

these three years I come seeking fruit on this fig tree, and find none: cut it down; why cumbereth it the ground?

dixit autem ad cultorem uineae ecce anni tres sunt ex quo uenio quaerens fructum in ficulnea hac et non inuenio succide ergo illam ut quid etiam terram occupat 13:8 at ille 13:8 And he answering said unto him, Lord, let it alone

this year also, till I shall dig about it, and dung it:

respondens dixit illi domine dimitte illam et hoc anno us-

que dum fodiam circa illam et mittam stercora 13:9 et si 13:9 And if it bear fruit, well: and if not, then after that

thou shalt cut it down.

quidem fecerit fructum sin autem in futurum succides eam 13:10 erat autem docens in synagoga eorum sabbatis 13:11 13:10 And he was teaching in one of the synagogues on

the sabbath.

13:11 And, behold, there was a woman which had a spirit of infirmity eighteen years, and was bowed together, and could in no wise lift up herself.

et ecce mulier quae habebat spiritum infirmitatis annis de-

Biblia Sacra Vulgata

cem et octo et erat inclinata nec omnino poterat sursum re-

13:12 And when Jesus saw her, he called her to him, and spicere 13:12 quam cum uideret Iesus uocauit ad se et ait

said unto her, Woman, thou art loosed from thine infirmity.

13:13 And he laid his hands on her: and immediately she illi mulier dimissa es ab infirmitate tua 13:13 et inposuit illi

was made straight, and glorified God.

13:14 And the ruler of the synagogue answered with in- manus et confestim erecta est et glorificabat Deum 13:14

dignation, because that Jesus had healed on the sabbath day, and said unto the people, There are six days in which men ought to work: in them therefore come and be healed, and not on the sabbath day.

respondens autem archisynagogus indignans quia sabbato curasset Iesus dicebat turbae sex dies sunt in quibus oportet operari in his ergo uenite et curamini et non in die sab-

13:15 The Lord then answered him, and said, Thou hyp- bati 13:15 respondit autem ad illum Dominus et dixit hypo-

ocrite, doth not each one of you on the sabbath loose his ox or his ass from the stall, and lead him away to watering?

critae unusquisque uestrum sabbato non soluit bouem suum

13:16 And ought not this woman, being a daughter of aut asinum a praesepio et ducit adaquare 13:16 hanc autem

Abraham, whom Satan hath bound, lo, these eighteen years, be loosed from this bond on the sabbath day?

filiam Abrahae quam alligauit Satanas ecce decem et octo

13:17 And when he had said these things, all his adver- annis non oportuit solui a uinculo isto die sabbati 13:17 et

saries were ashamed: and all the people rejoiced for all the glorious things that were done by him.

cum haec diceret erubescebant omnes aduersarii eius et omnis populus gaudebat in uniuersis quae gloriose fiebant ab eo

13:18 Then said he, Unto what is the kingdom of God like? 13:18 dicebat ergo cui simile est regnum Dei et cui simile

and whereunto shall I resemble it?

13:19 It is like a grain of mustard seed, which a man took, esse existimabo illud 13:19 simile est grano sinapis quod

and cast into his garden; and it grew, and waxed a great tree; and the fowls of the air lodged in the branches of it.

acceptum homo misit in hortum suum et creuit et factum est in arborem magnam et uolucres caeli requieuerunt in ra-

13:20 And again he said, Whereunto shall I liken the king- mis eius 13:20 et iterum dixit cui simile aestimabo regnum

dom of God?

13:21 It is like leaven, which a woman took and hid in Dei 13:21 simile est fermento quod acceptum mulier abs-

three measures of meal, till the whole was leavened.

13:22 And he went through the cities and villages, teach- condit in farinae sata tria donec fermentaretur totum 13:22

ing, and journeying toward Jerusalem.

et ibat per ciuitates et castella docens et iter faciens in Hieru-

13:23 Then said one unto him, Lord, are there few that be salem 13:23 ait autem illi quidam Domine si pauci sunt qui

saved? And he said unto them,

13:24 Strive to enter in at the strait gate: for many, I say saluantur ipse autem dixit ad illos 13:24 contendite intrare

unto you, will seek to enter in, and shall not be able.

per angustam portam quia multi dico uobis quaerunt intrare

13:25 When once the master of the house is risen up, and et non poterunt 13:25 cum autem intrauerit pater familias et

hath shut to the door, and ye begin to stand without, and to knock at the door, saying, Lord, Lord, open unto us; and he shall answer and say unto you, I know you not whence ye are:

cluserit ostium et incipietis foris stare et pulsare ostium dicentes Domine aperi nobis et respondens dicet uobis nescio

13:26 Then shall ye begin to say, We have eaten and drunk uos unde sitis 13:26 tunc incipietis dicere manducauimus

in thy presence, and thou hast taught in our streets.

13:27 But he shall say, I tell you, I know you not whence coram te et bibimus et in plateis nostris docuisti 13:27 et

ye are; depart from me, all ye workers of iniquity.

dicet uobis nescio uos unde sitis discedite a me omnes ope-

13:28 There shall be weeping and gnashing of teeth, when rarii iniquitatis 13:28 ibi erit fletus et stridor dentium cum

ye shall see Abraham, and Isaac, and Jacob, and all the prophets, in the kingdom of God, and you yourselves

thrust out. uideritis Abraham et Isaac et Iacob et omnes prophetas in re-

Euangelium Secundum Lucam

gno Dei uos autem expelli foras 13:29 et uenient ab oriente 13:29 And they shall come from the east, and from the

west, and from the north, and from the south, and shall sit down in the kingdom of God.

et occidente et aquilone et austro et accumbent in regno Dei 13:30 et ecce sunt nouissimi qui erunt primi et sunt primi qui 13:30 And, behold, there are last which shall be first, and

there are first which shall be last.

erunt nouissimi 13:31 in ipsa die accesserunt quidam Pha- 13:31 The same day there came certain of the Pharisees,

saying unto him, Get thee out, and depart hence: for Herod will kill thee.

risaeorum dicentes illi exi et uade hinc quia Herodes uult te occidere 13:32 et ait illis ite dicite uulpi illi ecce eicio 13:32 And he said unto them, Go ye, and tell that fox,

Behold, I cast out devils, and I do cures to day and to morrow, and the third day I shall be perfected.

daemonia et sanitates perficio hodie et cras et tertia consummor 13:33 uerumtamen oportet me hodie et cras et sequenti 13:33 Nevertheless I must walk to day, and to morrow, and

the day following: for it cannot be that a prophet perish out of Jerusalem.

ambulare quia non capit prophetam perire extra Hierusalem 13:34 Hierusalem Hierusalem quae occidis prophetas et lapi- 13:34 O Jerusalem, Jerusalem, which killest the prophets,

and stonest them that are sent unto thee; how often would I have gathered thy children together, as a hen doth gather

das eos qui mittuntur ad te quotiens uolui congregare filios her brood under her wings, and ye would not! tuos quemadmodum auis nidum suum sub pinnis et noluisti 13:35 ecce relinquitur uobis domus uestra dico autem uobis 13:35 Behold, your house is left unto you desolate: and

verily I say unto you, Ye shall not see me, until the time come when ye shall say, Blessed is he that cometh in the

quia non uidebitis me donec ueniat cum dicetis benedictus name of the Lord. qui uenit in nomine Domini

14:1 et factum est cum intraret in domum cuiusdam prin- 14:1 And it came to pass, as he went into the house of one

of the chief Pharisees to eat bread on the sabbath day, that they watched him.

cipis Pharisaeorum sabbato manducare panem et ipsi obseruabant eum 14:2 et ecce homo quidam hydropicus erat ante 14:2 And, behold, there was a certain man before him

which had the dropsy.

illum 14:3 et respondens Iesus dixit ad legis peritos et Pha- 14:3 And Jesus answering spake unto the lawyers and

Pharisees, saying, Is it lawful to heal on the sabbath day?

risaeos dicens si licet sabbato curare 14:4 at illi tacuerunt 14:4 And they held their peace. And he took him, and

healed him, and let him go;

ipse uero adprehensum sanauit eum ac dimisit 14:5 et res- 14:5 And answered them, saying, Which of you shall have

an ass or an ox fallen into a pit, and will not straightway pull him out on the sabbath day?

pondens ad illos dixit cuius uestrum asinus aut bos in puteum cadet et non continuo extrahet illum die sabbati 14:6 et non 14:6 And they could not answer him again to these things. poterant ad haec respondere illi 14:7 dicebat autem et ad in- 14:7 And he put forth a parable to those which were bid-

den, when he marked how they chose out the chief rooms; saying unto them.

uitatos parabolam intendens quomodo primos accubitus eligerent dicens ad illos 14:8 cum inuitatus fueris ad nuptias 14:8 When thou art bidden of any man to a wedding, sit

not down in the highest room; lest a more honourable man than thou be bidden of him;

non discumbas in primo loco ne forte honoratior te sit inuitatus ab eo 14:9 et ueniens is qui te et illum uocauit dicat 14:9 And he that bade thee and him come and say to thee,

Give this man place; and thou begin with shame to take the lowest room.

tibi da huic locum et tunc incipias cum rubore nouissimum locum tenere 14:10 sed cum uocatus fueris uade recumbe 14:10 But when thou art bidden, go and sit down in the

lowest room; that when he that bade thee cometh, he may say unto thee, Friend, go up higher: then shalt thou have

in nouissimo loco ut cum uenerit qui te inuitauit dicat tibi worship in the presence of them that sit at meat with thee. amice ascende superius tunc erit tibi gloria coram simul discumbentibus 14:11 quia omnis qui se exaltat humiliabitur et 14:11 For whosoever exalteth himself shall be abased; and

he that humbleth himself shall be exalted.

Biblia Sacra Vulgata

14:12 Then said he also to him that bade him, When thou qui se humiliat exaltabitur 14:12 dicebat autem et ei qui se

makest a dinner or a supper, call not thy friends, nor thy brethren, neither thy kinsmen, nor thy rich neighbours; lest

they also bid thee again, and a recompence be made thee. inuitauerat cum facis prandium aut cenam noli uocare ami-

cos tuos neque fratres tuos neque cognatos neque uicinos di-

14:13 But when thou makest a feast, call the poor, the uites ne forte et ipsi te reinuitent et fiat tibi retributio 14:13

maimed, the lame, the blind:

sed cum facis conuiuium uoca pauperes debiles claudos cae-

14:14 And thou shalt be blessed; for they cannot recom- cos 14:14 et beatus eris quia non habent retribuere tibi retri-

pense thee: for thou shalt be recompensed at the resurrection of the just.

buetur enim tibi in resurrectione iustorum 14:15 haec cum

14:15 And when one of them that sat at meat with him heard these things, he said unto him, Blessed is he that shall eat bread in the kingdom of God.

audisset quidam de simul discumbentibus dixit illi beatus qui

14:16 Then said he unto him, A certain man made a great manducabit panem in regno Dei 14:16 at ipse dixit ei homo

supper, and bade many:

14:17 And sent his servant at supper time to say to them quidam fecit cenam magnam et uocauit multos 14:17 et mi-

that were bidden, Come; for all things are now ready.

sit seruum suum hora cenae dicere inuitatis ut uenirent quia

14:18 And they all with one consent began to make ex- iam parata sunt omnia 14:18 et coeperunt simul omnes ex-

cuse. The first said unto him, I have bought a piece of ground, and I must needs go and see it: I pray thee have

me excused. cusare primus dixit ei uillam emi et necesse habeo exire et 14:19 And another said, I have bought five yoke of oxen, uidere illam rogo te habe me excusatum 14:19 et alter dixit

and I go to prove them: I pray thee have me excused.

iuga boum emi quinque et eo probare illa rogo te habe me

14:20 And another said, I have married a wife, and there- excusatum 14:20 et alius dixit uxorem duxi et ideo non pos-

fore I cannot come.

14:21 So that servant came, and shewed his lord these sum uenire 14:21 et reuersus seruus nuntiauit haec domino

things. Then the master of the house being angry said to his servant, Go out quickly into the streets and lanes of the city, and bring in hither the poor, and the maimed, and the halt, and the blind.

suo tunc iratus pater familias dixit seruo suo exi cito in plateas et uicos ciuitatis et pauperes ac debiles et caecos et clau-

14:22 And the servant said, Lord, it is done as thou hast dos introduc huc 14:22 et ait seruus domine factum est ut

commanded, and yet there is room.

14:23 And the lord said unto the servant, Go out into the imperasti et adhuc locus est 14:23 et ait dominus seruo exi

highways and hedges, and compel them to come in, that my house may be filled.

in uias et sepes et conpelle intrare ut impleatur domus mea

14:24 For I say unto you, That none of those men which 14:24 dico autem uobis quod nemo uirorum illorum qui uo-

were bidden shall taste of my supper.

14:25 And there went great multitudes with him: and he cati sunt gustabit cenam meam 14:25 ibant autem turbae

turned, and said unto them,

14:26 If any man come to me, and hate not his father, and multae cum eo et conuersus dixit ad illos 14:26 si quis ue-

mother, and wife, and children, and brethren, and sisters, yea, and his own life also, he cannot be my disciple.

nit ad me et non odit patrem suum et matrem et uxorem et filios et fratres et sorores adhuc autem et animam suam non

14:27 And whosoever doth not bear his cross, and come potest esse meus discipulus 14:27 et qui non baiulat cru-

after me, cannot be my disciple.

cem suam et uenit post me non potest esse meus discipu-

14:28 For which of you, intending to build a tower, sit- lus 14:28 quis enim ex uobis uolens turrem aedificare non

teth not down first, and counteth the cost, whether he have sufficient to finish it?

prius sedens conputat sumptus qui necessarii sunt si habet

14:29 Lest haply, after he hath laid the foundation, and is ad perficiendum 14:29 ne posteaquam posuerit fundamennot able to finish it, all that behold it begin to mock him,

tum et non potuerit perficere omnes qui uident incipiant in-

Euangelium Secundum Lucam

ludere ei 14:30 dicentes quia hic homo coepit aedificare et 14:30 Saying, This man began to build, and was not able

to finish.

non potuit consummare 14:31 aut qui rex iturus commit- 14:31 Or what king, going to make war against another

king, sitteth not down first, and consulteth whether he be able with ten thousand to meet him that cometh against

tere bellum aduersus alium regem non sedens prius cogitat him with twenty thousand? si possit cum decem milibus occurrere ei qui cum uiginti milibus uenit ad se 14:32 alioquin adhuc illo longe agente 14:32 Or else, while the other is yet a great way off, he

sendeth an ambassage, and desireth conditions of peace.

legationem mittens rogat ea quae pacis sunt 14:33 sic ergo 14:33 So likewise, whosoever he be of you that forsaketh

not all that he hath, he cannot be my disciple.

omnis ex uobis qui non renuntiat omnibus quae possidet non potest meus esse discipulus 14:34 bonum est sal si autem 14:34 Salt is good: but if the salt have lost his savour,

wherewith shall it be seasoned?

sal quoque euanuerit in quo condietur 14:35 neque in ter- 14:35 It is neither fit for the land, nor yet for the dunghill;

but men cast it out. He that hath ears to hear, let him hear.

ram neque in sterquilinium utile est sed foras mittetur qui habet aures audiendi audiat

15:1 erant autem adpropinquantes ei publicani et peccato- 15:1 Then drew near unto him all the publicans and sinners

for to hear him.

res ut audirent illum 15:2 et murmurabant Pharisaei et scri- 15:2 And the Pharisees and scribes murmured, saying,

This man receiveth sinners, and eateth with them.

bae dicentes quia hic peccatores recipit et manducat cum illis 15:3 et ait ad illos parabolam istam dicens 15:4 quis ex 15:3 And he spake this parable unto them, saying,

15:4 What man of you, having an hundred sheep, if he lose one of them, doth not leave the ninety and nine in the wilderness, and go after that which is lost, until he find it?

uobis homo qui habet centum oues et si perdiderit unam ex illis nonne dimittit nonaginta nouem in deserto et uadit ad illam quae perierat donec inueniat illam 15:5 et cum inuene- 15:5 And when he hath found it, he layeth it on his shoul-

ders, rejoicing.

rit eam inponit in umeros suos gaudens 15:6 et ueniens do- 15:6 And when he cometh home, he calleth together his

friends and neighbours, saying unto them, Rejoice with me; for I have found my sheep which was lost.

mum conuocat amicos et uicinos dicens illis congratulamini

mihi quia inueni ouem meam quae perierat 15:7 dico uo- 15:7 I say unto you, that likewise joy shall be in heaven over one sinner that repenteth, more than over ninety and nine just persons, which need no repentance.

bis quod ita gaudium erit in caelo super uno peccatore paenitentiam habente quam super nonaginta nouem iustis qui non indigent paenitentia 15:8 aut quae mulier habens drag- 15:8 Either what woman having ten pieces of silver, if

she lose one piece, doth not light a candle, and sweep the house, and seek diligently till she find it?

mas decem si perdiderit dragmam unam nonne accendit lucernam et euerrit domum et quaerit diligenter donec inueniat 15:9 et cum inuenerit conuocat amicas et uicinas di- 15:9 And when she hath found it, she calleth her friends

and her neighbours together, saying, Rejoice with me; for I have found the piece which I had lost.

cens congratulamini mihi quia inueni dragmam quam perdi-

deram 15:10 ita dico uobis gaudium erit coram angelis Dei 15:10 Likewise, I say unto you, there is joy in the presence

of the angels of God over one sinner that repenteth.

super uno peccatore paenitentiam agente 15:11 ait autem 15:11 And he said, A certain man had two sons: homo quidam habuit duos filios 15:12 et dixit adulescen- 15:12 And the younger of them said to his father, Father,

give me the portion of goods that falleth to me. And he divided unto them his living.

tior ex illis patri pater da mihi portionem substantiae quae me contingit et diuisit illis substantiam 15:13 et non post 15:13 And not many days after the younger son gathered

all together, and took his journey into a far country, and there wasted his substance with riotous living.

Biblia Sacra Vulgata

multos dies congregatis omnibus adulescentior filius peregre profectus est in regionem longinquam et ibi dissipauit sub-

15:14 And when he had spent all, there arose a mighty stantiam suam uiuendo luxuriose 15:14 et postquam omnia

famine in that land; and he began to be in want.

consummasset facta est fames ualida in regione illa et ipse

15:15 And he went and joined himself to a citizen of that coepit egere 15:15 et abiit et adhesit uni ciuium regionis il-

country; and he sent him into his fields to feed swine.

15:16 And he would fain have filled his belly with the lius et misit illum in uillam suam ut pasceret porcos 15:16 et

husks that the swine did eat: and no man gave unto him.

cupiebat implere uentrem suum de siliquis quas porci man-

15:17 And when he came to himself, he said, How many ducabant et nemo illi dabat 15:17 in se autem reuersus dixit

hired servants of my father’s have bread enough and to spare, and I perish with hunger!

quanti mercennarii patris mei abundant panibus ego autem

15:18 I will arise and go to my father, and will say unto hic fame pereo 15:18 surgam et ibo ad patrem meum et dihim, Father, I have sinned against heaven, and before thee,

15:19 And am no more worthy to be called thy son: make cam illi pater peccaui in caelum et coram te 15:19 et iam

me as one of thy hired servants.

non sum dignus uocari filius tuus fac me sicut unum de mer-

15:20 And he arose, and came to his father. But when cennariis tuis 15:20 et surgens uenit ad patrem suum cum

he was yet a great way off, his father saw him, and had compassion, and ran, and fell on his neck, and kissed him.

autem adhuc longe esset uidit illum pater ipsius et misericordia motus est et adcurrens cecidit supra collum eius et

15:21 And the son said unto him, Father, I have sinned osculatus est illum 15:21 dixitque ei filius pater peccaui in

against heaven, and in thy sight, and am no more worthy to be called thy son.

caelum et coram te iam non sum dignus uocari filius tuus

15:22 But the father said to his servants, Bring forth the 15:22 dixit autem pater ad seruos suos cito proferte stolam

best robe, and put it on him; and put a ring on his hand, and shoes on his feet:

primam et induite illum et date anulum in manum eius et

15:23 And bring hither the fatted calf, and kill it; and let calciamenta in pedes 15:23 et adducite uitulum saginatum us eat, and be merry:

15:24 For this my son was dead, and is alive again; he was et occidite et manducemus et epulemur 15:24 quia hic filius

lost, and is found. And they began to be merry.

meus mortuus erat et reuixit perierat et inuentus est et coe-

15:25 Now his elder son was in the field: and as he came perunt epulari 15:25 erat autem filius eius senior in agro et

and drew nigh to the house, he heard musick and dancing.

cum ueniret et adpropinquaret domui audiuit symphoniam

15:26 And he called one of the servants, and asked what et chorum 15:26 et uocauit unum de seruis et interrogauit

these things meant.

15:27 And he said unto him, Thy brother is come; and thy quae haec essent 15:27 isque dixit illi frater tuus uenit et

father hath killed the fatted calf, because he hath received him safe and sound.

occidit pater tuus uitulum saginatum quia saluum illum re-

15:28 And he was angry, and would not go in: therefore cepit 15:28 indignatus est autem et nolebat introire pater

came his father out, and intreated him.

15:29 And he answering said to his father, Lo, these many ergo illius egressus coepit rogare illum 15:29 at ille respon-

years do I serve thee, neither transgressed I at any time thy commandment: and yet thou never gavest me a kid, that I

might make merry with my friends: dens dixit patri suo ecce tot annis seruio tibi et numquam

mandatum tuum praeterii et numquam dedisti mihi hedum

15:30 But as soon as this thy son was come, which hath ut cum amicis meis epularer 15:30 sed postquam filius tuus

devoured thy living with harlots, thou hast killed for him the fatted calf.

hic qui deuorauit substantiam suam cum meretricibus uenit

Euangelium Secundum Lucam

occidisti illi uitulum saginatum 15:31 at ipse dixit illi fili tu 15:31 And he said unto him, Son, thou art ever with me,

and all that I have is thine.

semper mecum es et omnia mea tua sunt 15:32 epulari au- 15:32 It was meet that we should make merry, and be glad: for this thy brother was dead, and is alive again; and was lost, and is found.

tem et gaudere oportebat quia frater tuus hic mortuus erat et reuixit perierat et inuentus est

16:1 dicebat autem et ad discipulos suos homo quidam 16:1 And he said also unto his disciples, There was a certain rich man, which had a steward; and the same was accused unto him that he had wasted his goods.

erat diues qui habebat uilicum et hic diffamatus est apud illum quasi dissipasset bona ipsius 16:2 et uocauit illum et 16:2 And he called him, and said unto him, How is it that

I hear this of thee? give an account of thy stewardship; for thou mayest be no longer steward.

ait illi quid hoc audio de te redde rationem uilicationis tuae iam enim non poteris uilicare 16:3 ait autem uilicus intra 16:3 Then the steward said within himself, What shall I

do? for my lord taketh away from me the stewardship: I cannot dig; to beg I am ashamed.

se quid faciam quia dominus meus aufert a me uilicationem fodere non ualeo mendicare erubesco 16:4 scio quid 16:4 I am resolved what to do, that, when I am put out of

the stewardship, they may receive me into their houses.

faciam ut cum amotus fuero a uilicatione recipiant me in domos suas 16:5 conuocatis itaque singulis debitoribus do- 16:5 So he called every one of his lord’s debtors unto him,

and said unto the first, How much owest thou unto my lord?

mini sui dicebat primo quantum debes domino meo 16:6 at

16:6 And he said, An hundred measures of oil. And he said unto him, Take thy bill, and sit down quickly, and write fifty.

ille dixit centum cados olei dixitque illi accipe cautionem tuam et sede cito scribe quinquaginta 16:7 deinde alio dixit 16:7 Then said he to another, And how much owest thou?

And he said, An hundred measures of wheat. And he said unto him, Take thy bill, and write fourscore.

tu uero quantum debes qui ait centum choros tritici ait illi accipe litteras tuas et scribe octoginta 16:8 et laudauit do- 16:8 And the lord commended the unjust steward, because

he had done wisely: for the children of this world are in their generation wiser than the children of light.

minus uilicum iniquitatis quia prudenter fecisset quia filii huius saeculi prudentiores filiis lucis in generatione sua sunt 16:9 et ego uobis dico facite uobis amicos de mamona iniqui- 16:9 And I say unto you, Make to yourselves friends of

the mammon of unrighteousness; that, when ye fail, they may receive you into everlasting habitations.

tatis ut cum defeceritis recipiant uos in aeterna tabernacula 16:10 qui fidelis est in minimo et in maiori fidelis est et qui 16:10 He that is faithful in that which is least is faithful

also in much: and he that is unjust in the least is unjust also in much.

in modico iniquus est et in maiori iniquus est 16:11 si ergo

16:11 If therefore ye have not been faithful in the unrighteous mammon, who will commit to your trust the true riches?

in iniquo mamona fideles non fuistis quod uerum est quis credet uobis 16:12 et si in alieno fideles non fuistis quod 16:12 And if ye have not been faithful in that which is

another man’s, who shall give you that which is your own?

uestrum est quis dabit uobis 16:13 nemo seruus potest duo- 16:13 No servant can serve two masters: for either he will

hate the one, and love the other; or else he will hold to the one, and despise the other. Ye cannot serve God and

bus dominis seruire aut enim unum odiet et alterum diliget mammon. aut uni adherebit et alterum contemnet non potestis Deo seruire et mamonae 16:14 audiebant autem omnia haec Phari- 16:14 And the Pharisees also, who were covetous, heard

all these things: and they derided him.

saei qui erant auari et deridebant illum 16:15 et ait illis uos 16:15 And he said unto them, Ye are they which justify

yourselves before men; but God knoweth your hearts: for that which is highly esteemed among men is abomination

estis qui iustificatis uos coram hominibus Deus autem nouit in the sight of God. corda uestra quia quod hominibus altum est abominatio est

Biblia Sacra Vulgata

16:16 The law and the prophets were until John: since ante Deum 16:16 lex et prophetae usque ad Iohannem ex eo

that time the kingdom of God is preached, and every man presseth into it.

regnum Dei euangelizatur et omnis in illud uim facit 16:17

16:17 And it is easier for heaven and earth to pass, than one tittle of the law to fail.

facilius est autem caelum et terram praeterire quam de lege

16:18 Whosoever putteth away his wife, and marrieth an- unum apicem cadere 16:18 omnis qui dimittit uxorem suam other, committeth adultery: and whosoever marrieth her that is put away from her husband committeth adultery.

et ducit alteram moechatur et qui dimissam a uiro ducit moe-

16:19 There was a certain rich man, which was clothed in chatur 16:19 homo quidam erat diues et induebatur purpura

purple and fine linen, and fared sumptuously every day:

16:20 And there was a certain beggar named Lazarus, et bysso et epulabatur cotidie splendide 16:20 et erat qui-

which was laid at his gate, full of sores,

dam mendicus nomine Lazarus qui iacebat ad ianuam eius

16:21 And desiring to be fed with the crumbs which fell ulceribus plenus 16:21 cupiens saturari de micis quae ca-

from the rich man’s table: moreover the dogs came and licked his sores.

debant de mensa diuitis sed et canes ueniebant et lingebant

16:22 And it came to pass, that the beggar died, and was ulcera eius 16:22 factum est autem ut moreretur mendicus

carried by the angels into Abraham’s bosom: the rich man also died, and was buried;

et portaretur ab angelis in sinum Abrahae mortuus est autem

16:23 And in hell he lift up his eyes, being in torments, et diues et sepultus est in inferno 16:23 eleuans oculos suos

and seeth Abraham afar off, and Lazarus in his bosom.

cum esset in tormentis uidebat Abraham a longe et Laza-

16:24 And he cried and said, Father Abraham, have mercy rum in sinu eius 16:24 et ipse clamans dixit pater Abraham

on me, and send Lazarus, that he may dip the tip of his finger in water, and cool my tongue; for I am tormented in

this flame. miserere mei et mitte Lazarum ut intinguat extremum digiti

sui in aqua ut refrigeret linguam meam quia crucior in hac

16:25 But Abraham said, Son, remember that thou in thy flamma 16:25 et dixit illi Abraham fili recordare quia rece-

lifetime receivedst thy good things, and likewise Lazarus evil things: but now he is comforted, and thou art tor-

mented. pisti bona in uita tua et Lazarus similiter mala nunc autem 16:26 And beside all this, between us and you there is a hic consolatur tu uero cruciaris 16:26 et in his omnibus in-

great gulf fixed: so that they which would pass from hence to you cannot; neither can they pass to us, that would come

from thence. ter nos et uos chasma magnum firmatum est ut hii qui uolunt

hinc transire ad uos non possint neque inde huc transmeare

16:27 Then he said, I pray thee therefore, father, that thou 16:27 et ait rogo ergo te pater ut mittas eum in domum pawouldest send him to my father’s house:

16:28 For I have five brethren; that he may testify unto tris mei 16:28 habeo enim quinque fratres ut testetur illis

them, lest they also come into this place of torment.

16:29 Abraham saith unto him, They have Moses and the ne et ipsi ueniant in locum hunc tormentorum 16:29 et ait

prophets; let them hear them.

16:30 And he said, Nay, father Abraham: but if one went illi Abraham habent Mosen et prophetas audiant illos 16:30

unto them from the dead, they will repent.

at ille dixit non pater Abraham sed si quis ex mortuis ierit

16:31 And he said unto him, If they hear not Moses and ad eos paenitentiam agent 16:31 ait autem illi si Mosen et

the prophets, neither will they be persuaded, though one rose from the dead.

prophetas non audiunt neque si quis ex mortuis resurrexerit credent

17:1 Then said he unto the disciples, It is impossible but 17:1 et ad discipulos suos ait inpossibile est ut non ueniant

that offences will come: but woe unto him, through whom they come!

scandala uae autem illi per quem ueniunt 17:2 utilius est illi

17:2 It were better for him that a millstone were hanged about his neck, and he cast into the sea, than that he should offend one of these little ones.

si lapis molaris inponatur circa collum eius et proiciatur in mare quam ut scandalizet unum de pusillis istis 17:3 adten- 17:3 Take heed to yourselves: If thy brother trespass

against thee, rebuke him; and if he repent, forgive him.

dite uobis si peccauerit frater tuus increpa illum et si paenitentiam egerit dimitte illi 17:4 et si septies in die peccauerit 17:4 And if he trespass against thee seven times in a day,

and seven times in a day turn again to thee, saying, I repent; thou shalt forgive him.

in te et septies in die conuersus fuerit ad te dicens paenitet me dimitte illi 17:5 et dixerunt apostoli Domino adauge 17:5 And the apostles said unto the Lord, Increase our

faith.

nobis fidem 17:6 dixit autem Dominus si haberetis fidem 17:6 And the Lord said, If ye had faith as a grain of mus-

tard seed, ye might say unto this sycamine tree, Be thou plucked up by the root, and be thou planted in the sea; and

sicut granum sinapis diceretis huic arbori moro eradicare et it should obey you. transplantare in mare et oboediret uobis 17:7 quis autem 17:7 But which of you, having a servant plowing or feed-

ing cattle, will say unto him by and by, when he is come from the field, Go and sit down to meat?

uestrum habens seruum arantem aut pascentem qui regresso de agro dicet illi statim transi recumbe 17:8 et non dicet ei 17:8 And will not rather say unto him, Make ready where-

with I may sup, and gird thyself, and serve me, till I have eaten and drunken; and afterward thou shalt eat and drink?

para quod cenem et praecinge te et ministra mihi donec man-

ducem et bibam et post haec tu manducabis et bibes 17:9 17:9 Doth he thank that servant because he did the things

that were commanded him? I trow not.

numquid gratiam habet seruo illi quia fecit quae sibi imperauerat non puto 17:10 sic et uos cum feceritis omnia quae 17:10 So likewise ye, when ye shall have done all those

things which are commanded you, say, We are unprofitable servants: we have done that which was our duty to

praecepta sunt uobis dicite serui inutiles sumus quod debui- do. mus facere fecimus 17:11 et factum est dum iret in Hierusa- 17:11 And it came to pass, as he went to Jerusalem, that

he passed through the midst of Samaria and Galilee.

lem transiebat per mediam Samariam et Galilaeam 17:12 et 17:12 And as he entered into a certain village, there met

him ten men that were lepers, which stood afar off:

cum ingrederetur quoddam castellum occurrerunt ei decem uiri leprosi qui steterunt a longe 17:13 et leuauerunt uocem 17:13 And they lifted up their voices, and said, Jesus, Mas-

ter, have mercy on us.

dicentes Iesu praeceptor miserere nostri 17:14 quos ut uidit 17:14 And when he saw them, he said unto them, Go shew

yourselves unto the priests. And it came to pass, that, as they went, they were cleansed.

dixit ite ostendite uos sacerdotibus et factum est dum irent mundati sunt 17:15 unus autem ex illis ut uidit quia mun- 17:15 And one of them, when he saw that he was healed,

turned back, and with a loud voice glorified God,

datus est regressus est cum magna uoce magnificans Deum 17:16 et cecidit in faciem ante pedes eius gratias agens et hic 17:16 And fell down on his face at his feet, giving him

thanks: and he was a Samaritan.

erat Samaritanus 17:17 respondens autem Iesus dixit nonne 17:17 And Jesus answering said, Were there not ten

cleansed? but where are the nine?

decem mundati sunt et nouem ubi sunt 17:18 non est inuen- 17:18 There are not found that returned to give glory to

God, save this stranger.

tus qui rediiet et daret gloriam Deo nisi hic alienigena 17:19 17:19 And he said unto him, Arise, go thy way: thy faith

hath made thee whole.

et ait illi surge uade quia fides tua te saluum fecit 17:20 17:20 And when he was demanded of the Pharisees, when

the kingdom of God should come, he answered them and said, The kingdom of God cometh not with observation:

interrogatus autem a Pharisaeis quando uenit regnum Dei respondit eis et dixit non uenit regnum Dei cum obseruatione 17:21 neque dicent ecce hic aut ecce illic ecce enim 17:21 Neither shall they say, Lo here! or, lo there! for,

behold, the kingdom of God is within you.

regnum Dei intra uos est 17:22 et ait ad discipulos uenient 17:22 And he said unto the disciples, The days will come,

when ye shall desire to see one of the days of the Son of man, and ye shall not see it.

dies quando desideretis uidere unum diem Filii hominis et 17:23 And they shall say to you, See here; or, see there: non uidebitis 17:23 et dicent uobis ecce hic ecce illic no-

go not after them, nor follow them.

17:24 For as the lightning, that lighteneth out of the lite ire neque sectemini 17:24 nam sicut fulgur coruscans de

one part under heaven, shineth unto the other part under heaven; so shall also the Son of man be in his day.

sub caelo in ea quae sub caelo sunt fulget ita erit Filius homi-

17:25 But first must he suffer many things, and be rejected nis in die sua 17:25 primum autem oportet illum multa pati

of this generation.

17:26 And as it was in the days of Noe, so shall it be also et reprobari a generatione hac 17:26 et sicut factum est in

in the days of the Son of man.

17:27 They did eat, they drank, they married wives, they diebus Noe ita erit et in diebus Filii hominis 17:27 edebant

were given in marriage, until the day that Noe entered into the ark, and the flood came, and destroyed them all.

et bibebant uxores ducebant et dabantur ad nuptias usque in diem qua intrauit Noe in arcam et uenit diluuium et perdidit

17:28 Likewise also as it was in the days of Lot; they did omnes 17:28 similiter sicut factum est in diebus Loth ede-

eat, they drank, they bought, they sold, they planted, they builded;

bant et bibebant emebant et uendebant plantabant aedifica-

17:29 But the same day that Lot went out of Sodom it bant 17:29 qua die autem exiit Loth a Sodomis pluit ignem

rained fire and brimstone from heaven, and destroyed them all.

et sulphur de caelo et omnes perdidit 17:30 secundum haec

17:30 Even thus shall it be in the day when the Son of man is revealed.

17:31 In that day, he which shall be upon the housetop, erit qua die Filius hominis reuelabitur 17:31 in illa hora qui

and his stuff in the house, let him not come down to take it away: and he that is in the field, let him likewise not return

back. fuerit in tecto et uasa eius in domo ne descendat tollere illa 17:32 Remember Lot’s wife. et qui in agro similiter non redeet retro 17:32 memores es-

17:33 Whosoever shall seek to save his life shall lose it; tote uxoris Loth 17:33 quicumque quaesierit animam suam and whosoever shall lose his life shall preserve it.

saluare perdet illam et qui perdiderit illam uiuificabit eam

17:34 I tell you, in that night there shall be two men in one 17:34 dico uobis illa nocte erunt duo in lecto uno unus adbed; the one shall be taken, and the other shall be left.

17:35 Two women shall be grinding together; the one shall sumetur et alter relinquetur 17:35 duae erunt molentes in

be taken, and the other left.

unum una adsumetur et altera relinquetur duo in agro unus

17:36 Two men shall be in the field; the one shall be taken, adsumetur et alter relinquetur 17:36 respondentes dicunt illi

and the other left.

17:37 And they answered and said unto him, Where, ubi Domine 17:37 qui dixit eis ubicumque fuerit corpus il-

Lord? And he said unto them, Wheresoever the body is, thither will the eagles be gathered together.

luc congregabuntur aquilae

18:1 And he spake a parable unto them to this end, that 18:1 dicebat autem et parabolam ad illos quoniam oportet

men ought always to pray, and not to faint;

18:2 Saying, There was in a city a judge, which feared not semper orare et non deficere 18:2 dicens iudex quidam erat

God, neither regarded man:

in quadam ciuitate qui Deum non timebat et hominem non

18:3 And there was a widow in that city; and she came uerebatur 18:3 uidua autem quaedam erat in ciuitate illa et

unto him, saying, Avenge me of mine adversary.

18:4 And he would not for a while: but afterward he said ueniebat ad eum dicens uindica me de aduersario meo 18:4

within himself, Though I fear not God, nor regard man;

et nolebat per multum tempus post haec autem dixit intra se

18:5 Yet because this widow troubleth me, I will avenge et si Deum non timeo nec hominem reuereor 18:5 tamen her, lest by her continual coming she weary me.

quia molesta est mihi haec uidua uindicabo illam ne in no-

18:6 And the Lord said, Hear what the unjust judge saith. uissimo ueniens suggillet me 18:6 ait autem Dominus au-

18:7 And shall not God avenge his own elect, which cry dite quid iudex iniquitatis dicit 18:7 Deus autem non faciet

day and night unto him, though he bear long with them?

Euangelium Secundum Lucam

uindictam electorum suorum clamantium ad se die ac nocte et patientiam habebit in illis 18:8 dico uobis quia cito fa- 18:8 I tell you that he will avenge them speedily. Never-

theless when the Son of man cometh, shall he find faith on the earth?

ciet uindictam illorum uerumtamen Filius hominis ueniens

putas inueniet fidem in terra 18:9 dixit autem et ad quos- 18:9 And he spake this parable unto certain which trusted in themselves that they were righteous, and despised others:

dam qui in se confidebant tamquam iusti et aspernabantur ceteros parabolam istam 18:10 duo homines ascenderunt 18:10 Two men went up into the temple to pray; the one a

Pharisee, and the other a publican.

in templum ut orarent unus Pharisaeus et alter publicanus 18:11 Pharisaeus stans haec apud se orabat Deus gratias ago 18:11 The Pharisee stood and prayed thus with himself,

God, I thank thee, that I am not as other men are, extortioners, unjust, adulterers, or even as this publican.

tibi quia non sum sicut ceteri hominum raptores iniusti adulteri uel ut etiam hic publicanus 18:12 ieiuno bis in sabbato 18:12 I fast twice in the week, I give tithes of all that I

possess.

decimas do omnium quae possideo 18:13 et publicanus a 18:13 And the publican, standing afar off, would not lift up so much as his eyes unto heaven, but smote upon his breast, saying, God be merciful to me a sinner.

longe stans nolebat nec oculos ad caelum leuare sed percutiebat pectus suum dicens Deus propitius esto mihi peccatori 18:14 dico uobis descendit hic iustificatus in domum 18:14 I tell you, this man went down to his house justified

rather than the other: for every one that exalteth himself shall be abased; and he that humbleth himself shall be ex-

suam ab illo quia omnis qui se exaltat humiliabitur et qui alted. se humiliat exaltabitur 18:15 adferebant autem ad illum et 18:15 And they brought unto him also infants, that he

would touch them: but when his disciples saw it, they rebuked them.

infantes ut eos tangeret quod cum uiderent discipuli increpabant illos 18:16 Iesus autem conuocans illos dixit sinite 18:16 But Jesus called them unto him, and said, Suffer

little children to come unto me, and forbid them not: for of such is the kingdom of God.

pueros uenire ad me et nolite eos uetare talium est enim re-

gnum Dei 18:17 amen dico uobis quicumque non accepe- 18:17 Verily I say unto you, Whosoever shall not receive the kingdom of God as a little child shall in no wise enter therein.

rit regnum Dei sicut puer non intrabit in illud 18:18 et in-

18:18 And a certain ruler asked him, saying, Good Master, what shall I do to inherit eternal life?

terrogauit eum quidam princeps dicens magister bone quid faciens uitam aeternam possidebo 18:19 dixit autem ei Ie- 18:19 And Jesus said unto him, Why callest thou me

good? none is good, save one, that is, God.

sus quid me dicis bonum nemo bonus nisi solus Deus 18:20 18:20 Thou knowest the commandments, Do not commit

adultery, Do not kill, Do not steal, Do not bear false witness, Honour thy father and thy mother.

mandata nosti non occides non moechaberis non furtum facies non falsum testimonium dices honora patrem tuum et matrem 18:21 qui ait haec omnia custodiui a iuuentute mea 18:21 And he said, All these have I kept from my youth

up.

18:22 quo audito Iesus ait ei adhuc unum tibi deest omnia 18:22 Now when Jesus heard these things, he said unto

him, Yet lackest thou one thing: sell all that thou hast, and distribute unto the poor, and thou shalt have treasure in

quaecumque habes uende et da pauperibus et habebis the- heaven: and come, follow me. saurum in caelo et ueni sequere me 18:23 his ille auditis 18:23 And when he heard this, he was very sorrowful: for

he was very rich.

contristatus est quia diues erat ualde 18:24 uidens autem il- 18:24 And when Jesus saw that he was very sorrowful, he

said, How hardly shall they that have riches enter into the kingdom of God!

lum Iesus tristem factum dixit quam difficile qui pecunias habent in regnum Dei intrabunt 18:25 facilius est enim ca- 18:25 For it is easier for a camel to go through a needle’s

eye, than for a rich man to enter into the kingdom of God.

melum per foramen acus transire quam diuitem intrare in

18:26 And they that heard it said, Who then can be saved? regnum Dei 18:26 et dixerunt qui audiebant et quis potest 18:27 And he said, The things which are impossible with saluus fieri 18:27 ait illis quae inpossibilia sunt apud homi-

men are possible with God.

18:28 Then Peter said, Lo, we have left all, and followed nes possibilia sunt apud Deum 18:28 ait autem Petrus ecce

thee.

18:29 And he said unto them, Verily I say unto you, There nos dimisimus omnia et secuti sumus te 18:29 qui dixit eis

is no man that hath left house, or parents, or brethren, or wife, or children, for the kingdom of God’s sake,

amen dico uobis nemo est qui reliquit domum aut parentes

18:30 Who shall not receive manifold more in this present aut fratres aut uxorem aut filios propter regnum Dei 18:30 time, and in the world to come life everlasting.

et non recipiat multo plura in hoc tempore et in saeculo uen-

18:31 Then he took unto him the twelve, and said unto turo uitam aeternam 18:31 adsumpsit autem Iesus duode-

them, Behold, we go up to Jerusalem, and all things that are written by the prophets concerning the Son of man

shall be accomplished. cim et ait illis ecce ascendimus Hierosolyma et consummabuntur omnia quae scripta sunt per prophetas de Filio ho-

18:32 For he shall be delivered unto the Gentiles, and shall minis 18:32 tradetur enim gentibus et inludetur et flagelbe mocked, and spitefully entreated, and spitted on:

18:33 And they shall scourge him, and put him to death: labitur et conspuetur 18:33 et postquam flagellauerint oc-

and the third day he shall rise again.

18:34 And they understood none of these things: and this cident eum et die tertia resurget 18:34 et ipsi nihil horum

saying was hid from them, neither knew they the things which were spoken.

intellexerunt et erat uerbum istud absconditum ab eis et non

18:35 And it came to pass, that as he was come nigh unto intellegebant quae dicebantur 18:35 factum est autem cum

Jericho, a certain blind man sat by the way side begging:

adpropinquaret Hiericho caecus quidam sedebat secus uiam

18:36 And hearing the multitude pass by, he asked what it mendicans 18:36 et cum audiret turbam praetereuntem in-

meant.

18:37 And they told him, that Jesus of Nazareth passeth terrogabat quid hoc esset 18:37 dixerunt autem ei quod Ie-

by.

18:38 And he cried, saying, Jesus, thou son of David, have sus Nazarenus transiret 18:38 et clamauit dicens Iesu Fili

mercy on me.

18:39 And they which went before rebuked him, that he Dauid miserere mei 18:39 et qui praeibant increpabant eum

should hold his peace: but he cried so much the more, Thou son of David, have mercy on me.

ut taceret ipse uero multo magis clamabat Fili Dauid mise-

18:40 And Jesus stood, and commanded him to be brought rere mei 18:40 stans autem Iesus iussit illum adduci ad se et

unto him: and when he was come near, he asked him,

18:41 Saying, What wilt thou that I shall do unto thee? cum adpropinquasset interrogauit illum 18:41 dicens quid

And he said, Lord, that I may receive my sight.

18:42 And Jesus said unto him, Receive thy sight: thy faith tibi uis faciam at ille dixit Domine ut uideam 18:42 et Iesus

hath saved thee.

18:43 And immediately he received his sight, and fol- dixit illi respice fides tua te saluum fecit 18:43 et confestim

lowed him, glorifying God: and all the people, when they saw it, gave praise unto God.

uidit et sequebatur illum magnificans Deum et omnis plebs ut uidit dedit laudem Deo

19:1 And Jesus entered and passed through Jericho. 19:1 et ingressus perambulabat Hiericho 19:2 et ecce uir

19:2 And, behold, there was a man named Zacchaeus, which was the chief among the publicans, and he was rich.

nomine Zaccheus et hic erat princeps publicanorum et ipse

19:3 And he sought to see Jesus who he was; and could diues 19:3 et quaerebat uidere Iesum quis esset et non pote-

not for the press, because he was little of stature.

19:4 And he ran before, and climbed up into a sycomore rat prae turba quia statura pusillus erat 19:4 et praecurrens

tree to see him: for he was to pass that way.

ascendit in arborem sycomorum ut uideret illum quia inde

Euangelium Secundum Lucam

erat transiturus 19:5 et cum uenisset ad locum suspiciens 19:5 And when Jesus came to the place, he looked up, and

saw him, and said unto him, Zacchaeus, make haste, and come down; for to day I must abide at thy house.

Iesus uidit illum et dixit ad eum Zacchee festinans descende quia hodie in domo tua oportet me manere 19:6 et festi- 19:6 And he made haste, and came down, and received

him joyfully.

nans descendit et excepit illum gaudens 19:7 et cum uide- 19:7 And when they saw it, they all murmured, saying,

That he was gone to be guest with a man that is a sinner.

rent omnes murmurabant dicentes quod ad hominem peccatorem deuertisset 19:8 stans autem Zaccheus dixit ad Do- 19:8 And Zacchaeus stood, and said unto the Lord: Be-

hold, Lord, the half of my goods I give to the poor; and if I have taken any thing from any man by false accusation,

minum ecce dimidium bonorum meorum Domine do paupe- I restore him fourfold. ribus et si quid aliquem defraudaui reddo quadruplum 19:9 19:9 And Jesus said unto him, This day is salvation come

to this house, forsomuch as he also is a son of Abraham.

ait Iesus ad eum quia hodie salus domui huic facta est eo quod et ipse filius sit Abrahae 19:10 uenit enim Filius ho- 19:10 For the Son of man is come to seek and to save that

which was lost.

minis quaerere et saluum facere quod perierat 19:11 haec 19:11 And as they heard these things, he added and spake

a parable, because he was nigh to Jerusalem, and because they thought that the kingdom of God should immediately

illis audientibus adiciens dixit parabolam eo quod esset pro- appear. pe Hierusalem et quia existimarent quod confestim regnum Dei manifestaretur 19:12 dixit ergo homo quidam nobilis 19:12 He said therefore, A certain nobleman went into a

far country to receive for himself a kingdom, and to return.

abiit in regionem longinquam accipere sibi regnum et reuerti 19:13 uocatis autem decem seruis suis dedit illis de- 19:13 And he called his ten servants, and delivered them

ten pounds, and said unto them, Occupy till I come.

cem mnas et ait ad illos negotiamini dum uenio 19:14 ci- 19:14 But his citizens hated him, and sent a message after

him, saying, We will not have this man to reign over us.

ues autem eius oderant illum et miserunt legationem post illum dicentes nolumus hunc regnare super nos 19:15 et fac- 19:15 And it came to pass, that when he was returned,

having received the kingdom, then he commanded these servants to be called unto him, to whom he had given the money, that he might know how much every man had gained by trading.

tum est ut rediiet accepto regno et iussit uocari seruos quibus dedit pecuniam ut sciret quantum quisque negotiatus esset 19:16 uenit autem primus dicens domine mna tua de- 19:16 Then came the first, saying, Lord, thy pound hath

gained ten pounds.

cem mnas adquisiuit 19:17 et ait illi euge bone serue quia 19:17 And he said unto him, Well, thou good servant: be-

cause thou hast been faithful in a very little, have thou authority over ten cities.

in modico fidelis fuisti eris potestatem habens supra decem ciuitates 19:18 et alter uenit dicens domine mna tua fecit 19:18 And the second came, saying, Lord, thy pound hath

gained five pounds.

quinque mnas 19:19 et huic ait et tu esto supra quinque ciui- 19:19 And he said likewise to him, Be thou also over five

cities.

tates 19:20 et alter uenit dicens domine ecce mna tua quam 19:20 And another came, saying, Lord, behold, here is thy

pound, which I have kept laid up in a napkin:

habui repositam in sudario 19:21 timui enim te quia homo 19:21 For I feared thee, because thou art an austere man:

thou takest up that thou layedst not down, and reapest that thou didst not sow.

austeris es tollis quod non posuisti et metis quod non seminasti 19:22 dicit ei de ore tuo te iudico serue nequam scie- 19:22 And he saith unto him, Out of thine own mouth will

I judge thee, thou wicked servant. Thou knewest that I was an austere man, taking up that I laid not down, and reaping

bas quod ego austeris homo sum tollens quod non posui et that I did not sow: metens quod non seminaui 19:23 et quare non dedisti pe- 19:23 Wherefore then gavest not thou my money into the

bank, that at my coming I might have required mine own with usury?

cuniam meam ad mensam et ego ueniens cum usuris uti-

Biblia Sacra Vulgata

19:24 And he said unto them that stood by, Take from him que exegissem illud 19:24 et adstantibus dixit auferte ab illo

the pound, and give it to him that hath ten pounds. 19:25 (And they said unto him, Lord, he hath ten pounds.) mnam et date illi qui decem mnas habet 19:25 et dixerunt 19:26 For I say unto you, That unto every one which hath ei domine habet decem mnas 19:26 dico autem uobis quia shall be given; and from him that hath not, even that he hath shall be taken away from him.

omni habenti dabitur ab eo autem qui non habet et quod ha-

19:27 But those mine enemies, which would not that I bet auferetur ab eo 19:27 uerumtamen inimicos meos illos

should reign over them, bring hither, and slay them before me.

qui noluerunt me regnare super se adducite huc et interficite

19:28 And when he had thus spoken, he went before, as- ante me 19:28 et his dictis praecedebat ascendens in Hiero-

cending up to Jerusalem.

19:29 And it came to pass, when he was come nigh to solyma 19:29 et factum est cum adpropinquasset ad Beth-

Bethphage and Bethany, at the mount called the mount of Olives, he sent two of his disciples,

fage et Bethania ad montem qui uocatur Oliueti misit duos

19:30 Saying, Go ye into the village over against you; in discipulos suos 19:30 dicens ite in castellum quod contra

the which at your entering ye shall find a colt tied, whereon yet never man sat: loose him, and bring him hither.

est in quod introeuntes inuenietis pullum asinae alligatum cui nemo umquam hominum sedit soluite illum et adducite

19:31 And if any man ask you, Why do ye loose him? thus 19:31 et si quis uos interrogauerit quare soluitis sic dicetis ei

shall ye say unto him, Because the Lord hath need of him.

19:32 And they that were sent went their way, and found quia Dominus operam eius desiderat 19:32 abierunt autem

even as he had said unto them.

qui missi erant et inuenerunt sicut dixit illis stantem pullum

19:33 And as they were loosing the colt, the owners 19:33 soluentibus autem illis pullum dixerunt domini eius ad

thereof said unto them, Why loose ye the colt?

19:34 And they said, The Lord hath need of him. illos quid soluitis pullum 19:34 at illi dixerunt quia Domi-19:35 And they brought him to Jesus: and they cast their nus eum necessarium habet 19:35 et duxerunt illum ad Ie-

garments upon the colt, and they set Jesus thereon.

sum et iactantes uestimenta sua supra pullum inposuerunt Ie-

19:36 And as he went, they spread their clothes in the way. sum 19:36 eunte autem illo substernebant uestimenta sua in 19:37 And when he was come nigh, even now at the de- uia 19:37 et cum adpropinquaret iam ad descensum montis

scent of the mount of Olives, the whole multitude of the disciples began to rejoice and praise God with a loud voice

for all the mighty works that they had seen; Oliueti coeperunt omnes turbae discentium gaudentes lau-

dare Deum uoce magna super omnibus quas uiderant uirtuti-

19:38 Saying, Blessed be the King that cometh in the name bus 19:38 dicentes benedictus qui uenit rex in nomine Do-

of the Lord: peace in heaven, and glory in the highest.

19:39 And some of the Pharisees from among the multi- mini pax in caelo et gloria in excelsis 19:39 et quidam Pha-

tude said unto him, Master, rebuke thy disciples.

risaeorum de turbis dixerunt ad illum magister increpa disci-

19:40 And he answered and said unto them, I tell you that, pulos tuos 19:40 quibus ipse ait dico uobis quia si hii tacue-

if these should hold their peace, the stones would immediately cry out.

rint lapides clamabunt 19:41 et ut adpropinquauit uidens ci-

19:41 And when he was come near, he beheld the city, and wept over it,

19:42 Saying, If thou hadst known, even thou, at least in uitatem fleuit super illam dicens 19:42 quia si cognouisses

this thy day, the things which belong unto thy peace! but now they are hid from thine eyes.

et tu et quidem in hac die tua quae ad pacem tibi nunc au-

19:43 For the days shall come upon thee, that thine en- tem abscondita sunt ab oculis tuis 19:43 quia uenient dies in

emies shall cast a trench about thee, and compass thee round, and keep thee in on every side,

te et circumdabunt te inimici tui uallo et circumdabunt te et

19:44 And shall lay thee even with the ground, and thy coangustabunt te undique 19:44 ad terram prosternent te et

children within thee; and they shall not leave in thee one stone upon another; because thou knewest not the time of thy visitation.

Euangelium Secundum Lucam

filios qui in te sunt et non relinquent in te lapidem super lapidem eo quod non cognoueris tempus uisitationis tuae 19:45 19:45 And he went into the temple, and began to cast out

them that sold therein, and them that bought;

et ingressus in templum coepit eicere uendentes in illo et ementes 19:46 dicens illis scriptum est quia domus mea do- 19:46 Saying unto them, It is written, My house is the

house of prayer: but ye have made it a den of thieves.

mus orationis est uos autem fecistis illam speluncam latronum 19:47 et erat docens cotidie in templo principes autem 19:47 And he taught daily in the temple. But the chief

priests and the scribes and the chief of the people sought to destroy him,

sacerdotum et scribae et principes plebis quaerebant illum perdere 19:48 et non inueniebant quid facerent illi omnis 19:48 And could not find what they might do: for all the

people were very attentive to hear him.

enim populus suspensus erat audiens illum

20:1 et factum est in una dierum docente illo populum in 20:1 And it came to pass, that on one of those days, as he

taught the people in the temple, and preached the gospel, the chief priests and the scribes came upon him with the

templo et euangelizante conuenerunt principes sacerdotum elders, et scribae cum senioribus 20:2 et aiunt dicentes ad illum 20:2 And spake unto him, saying, Tell us, by what author-

ity doest thou these things? or who is he that gave thee this authority?

dic nobis in qua potestate haec facis aut quis est qui dedit tibi hanc potestatem 20:3 respondens autem dixit ad illos 20:3 And he answered and said unto them, I will also ask

you one thing; and answer me:

interrogabo uos et ego uerbum respondete mihi 20:4 bap- 20:4 The baptism of John, was it from heaven, or of men? tismum Iohannis de caelo erat an ex hominibus 20:5 at illi 20:5 And they reasoned with themselves, saying, If we

shall say, From heaven; he will say, Why then believed ye him not?

cogitabant inter se dicentes quia si dixerimus de caelo dicet quare ergo non credidistis illi 20:6 si autem dixerimus 20:6 But and if we say, Of men; all the people will stone

us: for they be persuaded that John was a prophet.

ex hominibus plebs uniuersa lapidabit nos certi sunt enim Iohannem prophetam esse 20:7 et responderunt se nescire 20:7 And they answered, that they could not tell whence it

was.

unde esset 20:8 et Iesus ait illis neque ego dico uobis in qua 20:8 And Jesus said unto them, Neither tell I you by what

authority I do these things.

potestate haec facio 20:9 coepit autem dicere ad plebem pa- 20:9 Then began he to speak to the people this parable; A

certain man planted a vineyard, and let it forth to husbandmen, and went into a far country for a long time.

rabolam hanc homo plantauit uineam et locauit eam colonis et ipse peregre fuit multis temporibus 20:10 et in tem- 20:10 And at the season he sent a servant to the husband-

men, that they should give him of the fruit of the vineyard: but the husbandmen beat him, and sent him away empty.

pore misit ad cultores seruum ut de fructu uineae darent illi qui caesum dimiserunt eum inanem 20:11 et addidit alte- 20:11 And again he sent another servant: and they beat

him also, and entreated him shamefully, and sent him away empty.

rum seruum mittere illi autem hunc quoque caedentes et adficientes contumelia dimiserunt inanem 20:12 et addidit ter- 20:12 And again he sent a third: and they wounded him

also, and cast him out.

tium mittere qui et illum uulnerantes eiecerunt 20:13 dixit 20:13 Then said the lord of the vineyard, What shall I do?

I will send my beloved son: it may be they will reverence him when they see him.

autem dominus uineae quid faciam mittam filium meum dilectum forsitan cum hunc uiderint uerebuntur 20:14 quem 20:14 But when the husbandmen saw him, they reasoned

among themselves, saying, This is the heir: come, let us kill him, that the inheritance may be ours.

cum uidissent coloni cogitauerunt inter se dicentes hic est heres occidamus illum ut nostra fiat hereditas 20:15 et eiec- 20:15 So they cast him out of the vineyard, and killed

him. What therefore shall the lord of the vineyard do unto them?

Biblia Sacra Vulgata

tum illum extra uineam occiderunt quid ergo faciet illis do-

20:16 He shall come and destroy these husbandmen, and minus uineae 20:16 ueniet et perdet colonos istos et dabit

shall give the vineyard to others. And when they heard it, they said, God forbid.

uineam aliis quo audito dixerunt illi absit 20:17 ille autem

20:17 And he beheld them, and said, What is this then that is written, The stone which the builders rejected, the same is become the head of the corner?

aspiciens eos ait quid est ergo hoc quod scriptum est lapidem quem reprobauerunt aedificantes hic factus est in ca-

20:18 Whosoever shall fall upon that stone shall be bro- put anguli 20:18 omnis qui ceciderit supra illum lapidem

ken; but on whomsoever it shall fall, it will grind him to powder.

conquassabitur supra quem autem ceciderit comminuet il-

20:19 And the chief priests and the scribes the same hour lum 20:19 et quaerebant principes sacerdotum et scribae

sought to lay hands on him; and they feared the people: for they perceived that he had spoken this parable against

them. mittere in illum manus illa hora et timuerunt populum co-

gnouerunt enim quod ad ipsos dixerit similitudinem istam

20:20 And they watched him, and sent forth spies, which 20:20 et obseruantes miserunt insidiatores qui se iustos simu-

should feign themselves just men, that they might take hold of his words, that so they might deliver him unto the

power and authority of the governor. larent ut caperent eum in sermone et traderent illum prin-20:21 And they asked him, saying, Master, we know that cipatui et potestati praesidis 20:21 et interrogauerunt illum

thou sayest and teachest rightly, neither acceptest thou the person of any, but teachest the way of God truly:

dicentes magister scimus quia recte dicis et doces et non

20:22 Is it lawful for us to give tribute unto Caesar, or no? accipis personam sed in ueritate uiam Dei doces 20:22 li-

20:23 But he perceived their craftiness, and said unto cet nobis dare tributum Caesari an non 20:23 considerans

them, Why tempt ye me?

20:24 Shew me a penny. Whose image and superscription autem dolum illorum dixit ad eos quid me temptatis 20:24

hath it? They answered and said, Caesar’s.

ostendite mihi denarium cuius habet imaginem et inscriptio-

20:25 And he said unto them, Render therefore unto Cae- nem respondentes dixerunt Caesaris 20:25 et ait illis reddite

sar the things which be Caesar’s, and unto God the things which be God’s.

ergo quae Caesaris sunt Caesari et quae Dei sunt Deo 20:26

20:26 And they could not take hold of his words before the people: and they marvelled at his answer, and held their peace.

et non potuerunt uerbum eius reprehendere coram plebe et

20:27 Then came to him certain of the Sadducees, which mirati in responso eius tacuerunt 20:27 accesserunt autem

deny that there is any resurrection; and they asked him,

quidam Sadducaeorum qui negant esse resurrectionem et in-

20:28 Saying, Master, Moses wrote unto us, If any man’s terrogauerunt eum 20:28 dicentes magister Moses scripsit

brother die, having a wife, and he die without children, that his brother should take his wife, and raise up seed

unto his brother. nobis si frater alicuius mortuus fuerit habens uxorem et hic

sine filiis fuerit ut accipiat eam frater eius uxorem et susci-

20:29 There were therefore seven brethren: and the first tet semen fratri suo 20:29 septem ergo fratres erant et pri-

took a wife, and died without children.

20:30 And the second took her to wife, and he died child- mus accepit uxorem et mortuus est sine filiis 20:30 et se-

less.

20:31 And the third took her; and in like manner the seven quens accepit illam et ipse mortuus est sine filio 20:31 et

also: and they left no children, and died.

tertius accepit illam similiter et omnes septem et non relique-

20:32 Last of all the woman died also. runt semen et mortui sunt 20:32 nouissima omnium mortua 20:33 Therefore in the resurrection whose wife of them is est et mulier 20:33 in resurrectione ergo cuius eorum erit

she? for seven had her to wife. 20:34 And Jesus answering said unto them, The children uxor siquidem septem habuerunt eam uxorem 20:34 et ait of this world marry, and are given in marriage:

Euangelium Secundum Lucam

illis Iesus filii saeculi huius nubunt et traduntur ad nuptias 20:35 illi autem qui digni habebuntur saeculo illo et resurrec- 20:35 But they which shall be accounted worthy to ob-

tain that world, and the resurrection from the dead, neither marry, nor are given in marriage:

tione ex mortuis neque nubunt neque ducunt uxores 20:36

20:36 Neither can they die any more: for they are equal unto the angels; and are the children of God, being the children of the resurrection.

neque enim ultra mori poterunt aequales enim angelis sunt et filii sunt Dei cum sint filii resurrectionis 20:37 quia uero 20:37 Now that the dead are raised, even Moses shewed at

the bush, when he calleth the Lord the God of Abraham, and the God of Isaac, and the God of Jacob.

resurgant mortui et Moses ostendit secus rubum sicut dicit Dominum Deum Abraham et Deum Isaac et Deum Iacob

20:38 Deus autem non est mortuorum sed uiuorum omnes 20:38 For he is not a God of the dead, but of the living: for

all live unto him.

enim uiuunt ei 20:39 respondentes autem quidam scriba- 20:39 Then certain of the scribes answering said, Master,

thou hast well said.

rum dixerunt magister bene dixisti 20:40 et amplius non 20:40 And after that they durst not ask him any question

at all.

audebant eum quicquam interrogare 20:41 dixit autem ad 20:41 And he said unto them, How say they that Christ is

David’s son?

illos quomodo dicunt Christum Filium Dauid esse 20:42 et 20:42 And David himself saith in the book of Psalms, The

LORD said unto my Lord, Sit thou on my right hand,

ipse Dauid dicit in libro Psalmorum dixit Dominus Domino meo sede a dextris meis 20:43 donec ponam inimicos tuos 20:43 Till I make thine enemies thy footstool.

scabillum pedum tuorum 20:44 Dauid ergo Dominum illum 20:44 David therefore calleth him Lord, how is he then his

son?

uocat et quomodo filius eius est 20:45 audiente autem omni 20:45 Then in the audience of all the people he said unto

his disciples,

populo dixit discipulis suis 20:46 adtendite a scribis qui uo- 20:46 Beware of the scribes, which desire to walk in long

robes, and love greetings in the markets, and the highest seats in the synagogues, and the chief rooms at feasts;

lunt ambulare in stolis et amant salutationes in foro et primas cathedras in synagogis et primos discubitus in conuiuiis 20:47 qui deuorant domos uiduarum simulantes longam ora- 20:47 Which devour widows’ houses, and for a shew make

long prayers: the same shall receive greater damnation.

tionem hii accipient damnationem maiorem

21:1 respiciens autem uidit eos qui mittebant munera sua 21:1 And he looked up, and saw the rich men casting their

gifts into the treasury.

in gazofilacium diuites 21:2 uidit autem et quandam ui- 21:2 And he saw also a certain poor widow casting in

thither two mites.

duam pauperculam mittentem aera minuta duo 21:3 et dixit 21:3 And he said, Of a truth I say unto you, that this poor

widow hath cast in more than they all:

uere dico uobis quia uidua haec pauper plus quam omnes

misit 21:4 nam omnes hii ex abundanti sibi miserunt in 21:4 For all these have of their abundance cast in unto the offerings of God: but she of her penury hath cast in all the living that she had.

munera Dei haec autem ex eo quod deest illi omnem uictum suum quem habuit misit 21:5 et quibusdam dicentibus 21:5 And as some spake of the temple, how it was adorned

with goodly stones and gifts, he said,

de templo quod lapidibus bonis et donis ornatum esset dixit 21:6 haec quae uidetis uenient dies in quibus non relinque- 21:6 As for these things which ye behold, the days will

come, in the which there shall not be left one stone upon another, that shall not be thrown down.

tur lapis super lapidem qui non destruatur 21:7 interroga-

21:7 And they asked him, saying, Master, but when shall these things be? and what sign will there be when these things shall come to pass?

uerunt autem illum dicentes praeceptor quando haec erunt et quod signum cum fieri incipient 21:8 qui dixit uidete ne 21:8 And he said, Take heed that ye be not deceived: for

many shall come in my name, saying, I am Christ; and the time draweth near: go ye not therefore after them.

Biblia Sacra Vulgata

seducamini multi enim uenient in nomine meo dicentes quia ego sum et tempus adpropinquauit nolite ergo ire post il-

21:9 But when ye shall hear of wars and commotions, be los 21:9 cum autem audieritis proelia et seditiones nolite

not terrified: for these things must first come to pass; but the end is not by and by.

terreri oportet primum haec fieri sed non statim finis 21:10

21:10 Then said he unto them, Nation shall rise against nation, and kingdom against kingdom:

tunc dicebat illis surget gens contra gentem et regnum ad-

21:11 And great earthquakes shall be in divers places, and uersus regnum 21:11 terraemotus magni erunt per loca et famines, and pestilences; and fearful sights and great signs shall there be from heaven.

pestilentiae et fames terroresque de caelo et signa magna

21:12 But before all these, they shall lay their hands on erunt 21:12 sed ante haec omnia inicient uobis manus suas

you, and persecute you, delivering you up to the synagogues, and into prisons, being brought before kings and

rulers for my name’s sake. et persequentur tradentes in synagogas et custodias trahen-21:13 And it shall turn to you for a testimony. tes ad reges et praesides propter nomen meum 21:13 con-

21:14 Settle it therefore in your hearts, not to meditate be- tinget autem uobis in testimonium 21:14 ponite ergo in corfore what ye shall answer:

dibus uestris non praemeditari quemadmodum respondeatis

21:15 For I will give you a mouth and wisdom, which all 21:15 ego enim dabo uobis os et sapientiam cui non pote-

your adversaries shall not be able to gainsay nor resist.

21:16 And ye shall be betrayed both by parents, and runt resistere et contradicere omnes aduersarii uestri 21:16

brethren, and kinsfolks, and friends; and some of you shall they cause to be put to death.

trademini autem a parentibus et fratribus et cognatis et ami-

21:17 And ye shall be hated of all men for my name’s sake. cis et morte adficient ex uobis 21:17 et eritis odio omnibus 21:18 But there shall not an hair of your head perish. propter nomen meum 21:18 et capillus de capite uestro non 21:19 In your patience possess ye your souls. peribit 21:19 in patientia uestra possidebitis animas uestras 21:20 And when ye shall see Jerusalem compassed with 21:20 cum autem uideritis circumdari ab exercitu Hierusa-

armies, then know that the desolation thereof is nigh.

21:21 Then let them which are in Judaea flee to the moun- lem tunc scitote quia adpropinquauit desolatio eius 21:21

tains; and let them which are in the midst of it depart out; and let not them that are in the countries enter thereinto.

tunc qui in Iudaea sunt fugiant in montes et qui in medio

21:22 For these be the days of vengeance, that all things eius discedant et qui in regionibus non intrent in eam 21:22

which are written may be fulfilled.

quia dies ultionis hii sunt ut impleantur omnia quae scripta

21:23 But woe unto them that are with child, and to them sunt 21:23 uae autem praegnatibus et nutrientibus in illis

that give suck, in those days! for there shall be great distress in the land, and wrath upon this people.

diebus erit enim pressura magna supra terram et ira populo

21:24 And they shall fall by the edge of the sword, and huic 21:24 et cadent in ore gladii et captiui ducentur in om-

shall be led away captive into all nations: and Jerusalem shall be trodden down of the Gentiles, until the times of

the Gentiles be fulfilled. nes gentes et Hierusalem calcabitur a gentibus donec imple-21:25 And there shall be signs in the sun, and in the moon, antur tempora nationum 21:25 et erunt signa in sole et luna

and in the stars; and upon the earth distress of nations, with perplexity; the sea and the waves roaring;

et stellis et in terris pressura gentium prae confusione soni-

21:26 Men’s hearts failing them for fear, and for looking tus maris et fluctuum 21:26 arescentibus hominibus prae ti-

after those things which are coming on the earth: for the powers of heaven shall be shaken.

more et expectatione quae superuenient uniuerso orbi nam

21:27 And then shall they see the Son of man coming in a uirtutes caelorum mouebuntur 21:27 et tunc uidebunt Fi-

cloud with power and great glory.

lium hominis uenientem in nube cum potestate magna et maiestate 21:28 his autem fieri incipientibus respicite et leuate 21:28 And when these things begin to come to pass, then

look up, and lift up your heads; for your redemption draweth nigh.

capita uestra quoniam adpropinquat redeeptio uestra 21:29

21:29 And he spake to them a parable; Behold the fig tree, and all the trees;

et dixit illis similitudinem uidete ficulneam et omnes arbores 21:30 cum producunt iam ex se fructum scitis quoniam pro- 21:30 When they now shoot forth, ye see and know of your

own selves that summer is now nigh at hand.

pe est aestas 21:31 ita et uos cum uideritis haec fieri scitote 21:31 So likewise ye, when ye see these things come to

pass, know ye that the kingdom of God is nigh at hand.

quoniam prope est regnum Dei 21:32 amen dico uobis quia 21:32 Verily I say unto you, This generation shall not pass

away, till all be fulfilled.

non praeteribit generatio haec donec omnia fiant 21:33 cae- 21:33 Heaven and earth shall pass away: but my words

shall not pass away.

lum et terra transibunt uerba autem mea non transient 21:34 21:34 And take heed to yourselves, lest at any time your

hearts be overcharged with surfeiting, and drunkenness, and cares of this life, and so that day come upon you un-

adtendite autem uobis ne forte grauentur corda uestra in cra- awares. pula et ebrietate et curis huius uitae et superueniat in uos repentina dies illa 21:35 tamquam laqueus enim superue- 21:35 For as a snare shall it come on all them that dwell

on the face of the whole earth.

niet in omnes qui sedent super faciem omnis terrae 21:36 21:36 Watch ye therefore, and pray always, that ye may

be accounted worthy to escape all these things that shall come to pass, and to stand before the Son of man.

uigilate itaque omni tempore orantes ut digni habeamini fugere ista omnia quae futura sunt et stare ante Filium hominis 21:37 erat autem diebus docens in templo noctibus uero ex- 21:37 And in the day time he was teaching in the temple;

and at night he went out, and abode in the mount that is called the mount of Olives.

iens morabatur in monte qui uocatur Oliueti 21:38 et omnis

21:38 And all the people came early in the morning to him in the temple, for to hear him.

populus manicabat ad eum in templo audire eum

22:1 adpropinquabat autem dies festus azymorum qui dici- 22:1 Now the feast of unleavened bread drew nigh, which

is called the Passover.

tur pascha 22:2 et quaerebant principes sacerdotum et scri- 22:2 And the chief priests and scribes sought how they

might kill him; for they feared the people.

bae quomodo eum interficerent timebant uero plebem 22:3 22:3 Then entered Satan into Judas surnamed Iscariot, be-

ing of the number of the twelve.

intrauit autem Satanas in Iudam qui cognominatur Scarioth unum de duodecim 22:4 et abiit et locutus est cum princi- 22:4 And he went his way, and communed with the chief

priests and captains, how he might betray him unto them.

pibus sacerdotum et magistratibus quemadmodum illum traderet eis 22:5 et gauisi sunt et pacti sunt pecuniam illi dare 22:5 And they were glad, and covenanted to give him

money.

22:6 et spopondit et quaerebat oportunitatem ut traderet il- 22:6 And he promised, and sought opportunity to betray

him unto them in the absence of the multitude.

lum sine turbis 22:7 uenit autem dies azymorum in qua ne- 22:7 Then came the day of unleavened bread, when the

passover must be killed.

cesse erat occidi pascha 22:8 et misit Petrum et Iohannem 22:8 And he sent Peter and John, saying, Go and prepare

us the passover, that we may eat.

dicens euntes parate nobis pascha ut manducemus 22:9 at 22:9 And they said unto him, Where wilt thou that we pre-

pare?

illi dixerunt ubi uis paremus 22:10 et dixit ad eos ecce in- 22:10 And he said unto them, Behold, when ye are entered

into the city, there shall a man meet you, bearing a pitcher of water; follow him into the house where he entereth in.

troeuntibus uobis in ciuitatem occurret uobis homo amphoram aquae portans sequimini eum in domum in qua intrat 22:11 et dicetis patri familias domus dicit tibi magister ubi 22:11 And ye shall say unto the goodman of the house,

The Master saith unto thee, Where is the guestchamber, where I shall eat the passover with my disciples?

est diuersorium ubi pascha cum discipulis meis manducem 22:12 And he shall shew you a large upper room fur- 22:12 et ipse uobis ostendet cenaculum magnum stratum et

nished: there make ready.

22:13 And they went, and found as he had said unto them: ibi parate 22:13 euntes autem inuenerunt sicut dixit illis et

and they made ready the passover.

22:14 And when the hour was come, he sat down, and the parauerunt pascha 22:14 et cum facta esset hora discubuit

twelve apostles with him.

22:15 And he said unto them, With desire I have desired et duodecim apostoli cum eo 22:15 et ait illis desiderio de-

to eat this passover with you before I suffer:

sideraui hoc pascha manducare uobiscum antequam patiar

22:16 For I say unto you, I will not any more eat thereof, 22:16 dico enim uobis quia ex hoc non manducabo illud do-

until it be fulfilled in the kingdom of God.

22:17 And he took the cup, and gave thanks, and said, nec impleatur in regno Dei 22:17 et accepto calice gratias

Take this, and divide it among yourselves:

22:18 For I say unto you, I will not drink of the fruit of the egit et dixit accipite et diuidite inter uos 22:18 dico enim

vine, until the kingdom of God shall come.

uobis quod non bibam de generatione uitis donec regnum

22:19 And he took bread, and gave thanks, and brake it, Dei ueniat 22:19 et accepto pane gratias egit et fregit et de-

and gave unto them, saying, This is my body which is given for you: this do in remembrance of me.

dit eis dicens hoc est corpus meum quod pro uobis datur

22:20 Likewise also the cup after supper, saying, This cup hoc facite in meam commemorationem 22:20 similiter et is the new testament in my blood, which is shed for you.

calicem postquam cenauit dicens hic est calix nouum tes-

22:21 But, behold, the hand of him that betrayeth me is tamentum in sanguine meo quod pro uobis funditur 22:21

with me on the table.

uerumtamen ecce manus tradentis me mecum est in mensa

22:22 And truly the Son of man goeth, as it was deter- 22:22 et quidem Filius hominis secundum quod definitum est

mined: but woe unto that man by whom he is betrayed!

22:23 And they began to enquire among themselves, uadit uerumtamen uae illi homini per quem traditur 22:23

which of them it was that should do this thing.

et ipsi coeperunt quaerere inter se quis esset ex eis qui hoc

22:24 And there was also a strife among them, which of facturus esset 22:24 facta est autem et contentio inter eos

them should be accounted the greatest.

22:25 And he said unto them, The kings of the Gentiles quis eorum uideretur esse maior 22:25 dixit autem eis re-

exercise lordship over them; and they that exercise authority upon them are called benefactors.

ges gentium dominantur eorum et qui potestatem habent su-

22:26 But ye shall not be so: but he that is greatest among per eos benefici uocantur 22:26 uos autem non sic sed qui you, let him be as the younger; and he that is chief, as he that doth serve.

maior est in uobis fiat sicut iunior et qui praecessor est sicut

22:27 For whether is greater, he that sitteth at meat, or he ministrator 22:27 nam quis maior est qui recumbit an qui that serveth? is not he that sitteth at meat? but I am among you as he that serveth.

ministrat nonne qui recumbit ego autem in medio uestrum

22:28 Ye are they which have continued with me in my sum sicut qui ministrat 22:28 uos autem estis qui perman-

temptations.

22:29 And I appoint unto you a kingdom, as my Father sistis mecum in temptationibus meis 22:29 et ego dispono

hath appointed unto me;

22:30 That ye may eat and drink at my table in my king- uobis sicut disposuit mihi Pater meus regnum 22:30 ut eda-

dom, and sit on thrones judging the twelve tribes of Israel.

tis et bibatis super mensam meam in regno et sedeatis super

22:31 And the Lord said, Simon, Simon, behold, Satan thronos iudicantes duodecim tribus Israhel 22:31 ait autem

hath desired to have you, that he may sift you as wheat:

Dominus Simon Simon ecce Satanas expetiuit uos ut crib-

22:32 But I have prayed for thee, that thy faith fail not: raret sicut triticum 22:32 ego autem rogaui pro te ut non and when thou art converted, strengthen thy brethren.

deficiat fides tua et tu aliquando conuersus confirma fratres tuos 22:33 qui dixit ei Domine tecum paratus sum et in car- 22:33 And he said unto him, Lord, I am ready to go with

thee, both into prison, and to death.

cerem et in mortem ire 22:34 et ille dixit dico tibi Petre non 22:34 And he said, I tell thee, Peter, the cock shall not

crow this day, before that thou shalt thrice deny that thou knowest me.

cantabit hodie gallus donec ter abneges nosse me 22:35 et

22:35 And he said unto them, When I sent you without purse, and scrip, and shoes, lacked ye any thing? And they said, Nothing.

dixit eis quando misi uos sine sacculo et pera et calciamentis numquid aliquid defuit uobis at illi dixerunt nihil 22:36 22:36 Then said he unto them, But now, he that hath a

purse, let him take it, and likewise his scrip: and he that hath no sword, let him sell his garment, and buy one.

dixit ergo eis sed nunc qui habet sacculum tollat similiter et peram et qui non habet uendat tunicam suam et emat gla-

dium 22:37 dico enim uobis quoniam adhuc hoc quod scri- 22:37 For I say unto you, that this that is written must yet be accomplished in me, And he was reckoned among the transgressors: for the things concerning me have an end.

ptum est oportet impleri in me et quod cum iniustis deputatus est etenim ea quae sunt de me finem habent 22:38 at illi 22:38 And they said, Lord, behold, here are two swords.

And he said unto them, It is enough.

dixerunt Domine ecce gladii duo hic at ille dixit eis satis est 22:39 et egressus ibat secundum consuetudinem in montem 22:39 And he came out, and went, as he was wont, to the

mount of Olives; and his disciples also followed him.

Oliuarum secuti sunt autem illum et discipuli 22:40 et cum 22:40 And when he was at the place, he said unto them,

Pray that ye enter not into temptation.

peruenisset ad locum dixit illis orate ne intretis in temptationem 22:41 et ipse auulsus est ab eis quantum iactus est 22:41 And he was withdrawn from them about a stone’s

cast, and kneeled down, and prayed,

lapidis et positis genibus orabat 22:42 dicens Pater si uis 22:42 Saying, Father, if thou be willing, remove this cup

from me: nevertheless not my will, but thine, be done.

transfer calicem istum a me uerumtamen non mea uoluntas sed tua fiat 22:43 apparuit autem illi angelus de caelo 22:43 And there appeared an angel unto him from heaven,

strengthening him.

confortans eum et factus in agonia prolixius orabat 22:44 22:44 And being in an agony he prayed more earnestly:

and his sweat was as it were great drops of blood falling down to the ground.

et factus est sudor eius sicut guttae sanguinis decurrentis in terram 22:45 et cum surrexisset ab oratione et uenisset ad 22:45 And when he rose up from prayer, and was come to

his disciples, he found them sleeping for sorrow,

discipulos suos inuenit eos dormientes prae tristitia 22:46 22:46 And said unto them, Why sleep ye? rise and pray,

lest ye enter into temptation.

et ait illis quid dormitis surgite orate ne intretis in temptationem 22:47 adhuc eo loquente ecce turba et qui uocaba- 22:47 And while he yet spake, behold a multitude, and he

that was called Judas, one of the twelve, went before them, and drew near unto Jesus to kiss him.

tur Iudas unus de duodecim antecedebat eos et adpropinquauit Iesu ut oscularetur eum 22:48 Iesus autem dixit ei Iuda 22:48 But Jesus said unto him, Judas, betrayest thou the

Son of man with a kiss?

osculo Filium hominis tradis 22:49 uidentes autem hii qui 22:49 When they which were about him saw what would

follow, they said unto him, Lord, shall we smite with the sword?

circa ipsum erant quod futurum erat dixerunt ei Domine si percutimus in gladio 22:50 et percussit unus ex illis seruum 22:50 And one of them smote the servant of the high

priest, and cut off his right ear.

principis sacerdotum et amputauit auriculam eius dextram 22:51 respondens autem Iesus ait sinite usque huc et cum te- 22:51 And Jesus answered and said, Suffer ye thus far.

And he touched his ear, and healed him.

tigisset auriculam eius sanauit eum 22:52 dixit autem Iesus 22:52 Then Jesus said unto the chief priests, and captains

of the temple, and the elders, which were come to him, Be ye come out, as against a thief, with swords and staves?

ad eos qui uenerant ad se principes sacerdotum et magistra-

Biblia Sacra Vulgata

tus templi et seniores quasi ad latronem existis cum gladiis et

22:53 When I was daily with you in the temple, ye fustibus 22:53 cum cotidie uobiscum fuerim in templo non

stretched forth no hands against me: but this is your hour, and the power of darkness.

extendistis manus in me sed haec est hora uestra et potestas

22:54 Then took they him, and led him, and brought him tenebrarum 22:54 conprehendentes autem eum duxerunt ad

into the high priest’s house. And Peter followed afar off.

domum principis sacerdotum Petrus uero sequebatur a longe

22:55 And when they had kindled a fire in the midst of the 22:55 accenso autem igni in medio atrio et circumsedentibus

hall, and were set down together, Peter sat down among them.

illis erat Petrus in medio eorum 22:56 quem cum uidisset

22:56 But a certain maid beheld him as he sat by the fire, and earnestly looked upon him, and said, This man was also with him.

ancilla quaedam sedentem ad lumen et eum fuisset intuita

22:57 And he denied him, saying, Woman, I know him dixit et hic cum illo erat 22:57 at ille negauit eum dicens

not.

22:58 And after a little while another saw him, and said, mulier non noui illum 22:58 et post pusillum alius uidens

Thou art also of them. And Peter said, Man, I am not.

eum dixit et tu de illis es Petrus uero ait o homo non sum

22:59 And about the space of one hour after another con- 22:59 et interuallo facto quasi horae unius alius quidam ad-

fidently affirmed, saying, Of a truth this fellow also was with him: for he is a Galilaean.

firmabat dicens uere et hic cum illo erat nam et Galilaeus est

22:60 And Peter said, Man, I know not what thou sayest. 22:60 et ait Petrus homo nescio quod dicis et continuo adhuc

And immediately, while he yet spake, the cock crew.

22:61 And the Lord turned, and looked upon Peter. And illo loquente cantauit gallus 22:61 et conuersus Dominus

Peter remembered the word of the Lord, how he had said unto him, Before the cock crow, thou shalt deny me thrice.

respexit Petrum et recordatus est Petrus uerbi Domini sicut

22:62 And Peter went out, and wept bitterly. dixit quia priusquam gallus cantet ter me negabis 22:62 et 22:63 And the men that held Jesus mocked him, and smote egressus foras Petrus fleuit amare 22:63 et uiri qui tenebant

him.

22:64 And when they had blindfolded him, they struck illum inludebant ei caedentes 22:64 et uelauerunt eum et

him on the face, and asked him, saying, Prophesy, who is it that smote thee?

percutiebant faciem eius et interrogabant eum dicentes pro-

22:65 And many other things blasphemously spake they phetiza quis est qui te percussit 22:65 et alia multa blas-

against him.

22:66 And as soon as it was day, the elders of the people phemantes dicebant in eum 22:66 et ut factus est dies con-

and the chief priests and the scribes came together, and led him into their council, saying,

uenerunt seniores plebis et principes sacerdotum et scribae et duxerunt illum in concilium suum dicentes si tu es Chri-

22:67 Art thou the Christ? tell us. And he said unto them, stus dic nobis 22:67 et ait illis si uobis dixero non creditis

If I tell you, ye will not believe:

22:68 And if I also ask you, ye will not answer me, nor let mihi 22:68 si autem et interrogauero non respondebitis mihi

me go.

22:69 Hereafter shall the Son of man sit on the right hand neque dimittetis 22:69 ex hoc autem erit Filius hominis se-

of the power of God.

22:70 Then said they all, Art thou then the Son of God? dens a dextris uirtutis Dei 22:70 dixerunt autem omnes tu

And he said unto them, Ye say that I am.

22:71 And they said, What need we any further witness? ergo es Filius Dei qui ait uos dicitis quia ego sum 22:71 at

for we ourselves have heard of his own mouth.

illi dixerunt quid adhuc desideramus testimonium ipsi enim audiuimus de ore eius

23:1 And the whole multitude of them arose, and led him 23:1 et surgens omnis multitudo eorum duxerunt illum ad

unto Pilate.

23:2 And they began to accuse him, saying, We found this Pilatum 23:2 coeperunt autem accusare illum dicentes hunc

fellow perverting the nation, and forbidding to give tribute to Caesar, saying that he himself is Christ a King.

Euangelium Secundum Lucam

inuenimus subuertentem gentem nostram et prohibentem tributa dari Caesari et dicentem se Christum regem esse 23:3 23:3 And Pilate asked him, saying, Art thou the King of

the Jews? And he answered him and said, Thou sayest it.

Pilatus autem interrogauit eum dicens tu es rex Iudaeorum at ille respondens ait tu dicis 23:4 ait autem Pilatus ad prin- 23:4 Then said Pilate to the chief priests and to the people,

I find no fault in this man.

cipes sacerdotum et turbas nihil inuenio causae in hoc homine 23:5 at illi inualescebant dicentes commouet popu- 23:5 And they were the more fierce, saying, He stirreth up

the people, teaching throughout all Jewry, beginning from Galilee to this place.

lum docens per uniuersam Iudaeam et incipiens a Galilaea usque huc 23:6 Pilatus autem audiens Galilaeam interro- 23:6 When Pilate heard of Galilee, he asked whether the

man were a Galilaean.

gauit si homo Galilaeus esset 23:7 et ut cognouit quod de 23:7 And as soon as he knew that he belonged unto

Herod’s jurisdiction, he sent him to Herod, who himself also was at Jerusalem at that time.

Herodis potestate esset remisit eum ad Herodem qui et ipse Hierosolymis erat illis diebus 23:8 Herodes autem uiso Iesu 23:8 And when Herod saw Jesus, he was exceeding glad:

for he was desirous to see him of a long season, because he had heard many things of him; and he hoped to have

gauisus est ualde erat enim cupiens ex multo tempore uidere seen some miracle done by him. eum eo quod audiret multa de illo et sperabat signum aliquod uidere ab eo fieri 23:9 interrogabat autem illum multis ser- 23:9 Then he questioned with him in many words; but he

answered him nothing.

monibus at ipse nihil illi respondebat 23:10 stabant etiam 23:10 And the chief priests and scribes stood and vehe-

mently accused him.

principes sacerdotum et scribae constanter accusantes eum 23:11 spreuit autem illum Herodes cum exercitu suo et inlu- 23:11 And Herod with his men of war set him at nought,

and mocked him, and arrayed him in a gorgeous robe, and sent him again to Pilate.

sit indutum ueste alba et remisit ad Pilatum 23:12 et facti

23:12 And the same day Pilate and Herod were made friends together: for before they were at enmity between themselves.

sunt amici Herodes et Pilatus in ipsa die nam antea inimici erant ad inuicem 23:13 Pilatus autem conuocatis principi- 23:13 And Pilate, when he had called together the chief

priests and the rulers and the people,

bus sacerdotum et magistratibus et plebe 23:14 dixit ad il- 23:14 Said unto them, Ye have brought this man unto me,

as one that perverteth the people: and, behold, I, having examined him before you, have found no fault in this man

los obtulistis mihi hunc hominem quasi auertentem populum touching those things whereof ye accuse him: et ecce ego coram uobis interrogans nullam causam inueni in homine isto ex his in quibus eum accusatis 23:15 sed ne- 23:15 No, nor yet Herod: for I sent you to him; and, lo,

nothing worthy of death is done unto him.

que Herodes nam remisi uos ad illum et ecce nihil dignum morte actum est ei 23:16 emendatum ergo illum dimittam 23:16 I will therefore chastise him, and release him. 23:17 necesse autem habebat dimittere eis per diem festum 23:17 (For of necessity he must release one unto them at

the feast.)

unum 23:18 exclamauit autem simul uniuersa turba dicens 23:18 And they cried out all at once, saying, Away with

this man, and release unto us Barabbas:

tolle hunc et dimitte nobis Barabban 23:19 qui erat propter 23:19 (Who for a certain sedition made in the city, and for

murder, was cast into prison.)

seditionem quandam factam in ciuitate et homicidium missus in carcerem 23:20 iterum autem Pilatus locutus est ad 23:20 Pilate therefore, willing to release Jesus, spake

again to them.

illos uolens dimittere Iesum 23:21 at illi succlamabant di- 23:21 But they cried, saying, Crucify him, crucify him. centes crucifige crucifige illum 23:22 ille autem tertio dixit 23:22 And he said unto them the third time, Why, what

evil hath he done? I have found no cause of death in him: I will therefore chastise him, and let him go.

Biblia Sacra Vulgata

ad illos quid enim mali fecit iste nullam causam mortis in-

23:23 And they were instant with loud voices, requiring uenio in eo corripiam ergo illum et dimittam 23:23 at illi

that he might be crucified. And the voices of them and of the chief priests prevailed.

instabant uocibus magnis postulantes ut crucifigeretur et in-

23:24 And Pilate gave sentence that it should be as they ualescebant uoces eorum 23:24 et Pilatus adiudicauit fieri

required. 23:25 And he released unto them him that for sedition and petitionem eorum 23:25 dimisit autem illis eum qui promurder was cast into prison, whom they had desired; but he delivered Jesus to their will.

pter homicidium et seditionem missus fuerat in carcerem

23:26 And as they led him away, they laid hold upon one quem petebant Iesum uero tradidit uoluntati eorum 23:26 et

Simon, a Cyrenian, coming out of the country, and on him they laid the cross, that he might bear it after Jesus.

cum ducerent eum adprehenderunt Simonem quendam Cyrenensem uenientem de uilla et inposuerunt illi crucem por-

23:27 And there followed him a great company of people, tare post Iesum 23:27 sequebatur autem illum multa turba

and of women, which also bewailed and lamented him.

populi et mulierum quae plangebant et lamentabant eum

23:28 But Jesus turning unto them said, Daughters of 23:28 conuersus autem ad illas Iesus dixit filiae Hierusalem

Jerusalem, weep not for me, but weep for yourselves, and for your children.

nolite flere super me sed super uos ipsas flete et super fi-

23:29 For, behold, the days are coming, in the which they lios uestros 23:29 quoniam ecce uenient dies in quibus di-

shall say, Blessed are the barren, and the wombs that never bare, and the paps which never gave suck.

cent beatae steriles et uentres qui non genuerunt et ubera

23:30 Then shall they begin to say to the mountains, Fall quae non lactauerunt 23:30 tunc incipient dicere montibus

on us; and to the hills, Cover us.

23:31 For if they do these things in a green tree, what shall cadite super nos et collibus operite nos 23:31 quia si in ui-

be done in the dry? 23:32 And there were also two other, malefactors, led with ridi ligno haec faciunt in arido quid fiet 23:32 ducebantur him to be put to death.

23:33 And when they were come to the place, which is autem et alii duo nequam cum eo ut interficerentur 23:33 et

called Calvary, there they crucified him, and the malefactors, one on the right hand, and the other on the left.

postquam uenerunt in locum qui uocatur Caluariae ibi crucifixerunt eum et latrones unum a dextris et alterum a sinistris

23:34 Then said Jesus, Father, forgive them; for they know 23:34 Iesus autem dicebat Pater dimitte illis non enim sciunt

not what they do. And they parted his raiment, and cast lots.

quid faciunt diuidentes uero uestimenta eius miserunt sortes

23:35 And the people stood beholding. And the rulers also 23:35 et stabat populus expectans et deridebant illum prin-

with them derided him, saying, He saved others; let him save himself, if he be Christ, the chosen of God.

cipes cum eis dicentes alios saluos fecit se saluum faciat si

23:36 And the soldiers also mocked him, coming to him, hic est Christus Dei electus 23:36 inludebant autem ei et

and offering him vinegar,

23:37 And saying, If thou be the king of the Jews, save milites accedentes et acetum offerentes illi 23:37 dicentes

thyself. 23:38 And a superscription also was written over him in si tu es rex Iudaeorum saluum te fac 23:38 erat autem et letters of Greek, and Latin, and Hebrew, THIS IS THE KING OF THE JEWS.

superscriptio inscripta super illum litteris graecis et latinis

23:39 And one of the malefactors which were hanged et hebraicis hic est rex Iudaeorum 23:39 unus autem de his

railed on him, saying, If thou be Christ, save thyself and us.

qui pendebant latronibus blasphemabat eum dicens si tu es

23:40 But the other answering rebuked him, saying, Dost Christus saluum fac temet ipsum et nos 23:40 respondens not thou fear God, seeing thou art in the same condemnation?

autem alter increpabat illum dicens neque tu times Deum

Euangelium Secundum Lucam

quod in eadem damnatione es 23:41 et nos quidem iuste 23:41 And we indeed justly; for we receive the due reward

of our deeds: but this man hath done nothing amiss.

nam digna factis recipimus hic uero nihil mali gessit 23:42 23:42 And he said unto Jesus, Lord, remember me when

thou comest into thy kingdom.

et dicebat ad Iesum Domine memento mei cum ueneris in regnum tuum 23:43 et dixit illi Iesus amen dico tibi hodie 23:43 And Jesus said unto him, Verily I say unto thee, To

day shalt thou be with me in paradise.

mecum eris in paradiso 23:44 erat autem fere hora sexta et 23:44 And it was about the sixth hour, and there was a

darkness over all the earth until the ninth hour.

tenebrae factae sunt in uniuersa terra usque in nonam horam 23:45 et obscuratus est sol et uelum templi scissum est 23:45 And the sun was darkened, and the veil of the temple

was rent in the midst.

medium 23:46 et clamans uoce magna Iesus ait Pater in ma- 23:46 And when Jesus had cried with a loud voice, he said,

Father, into thy hands I commend my spirit: and having said thus, he gave up the ghost.

nus tuas commendo spiritum meum et haec dicens exspirauit 23:47 uidens autem centurio quod factum fuerat glorifi- 23:47 Now when the centurion saw what was done, he

glorified God, saying, Certainly this was a righteous man.

cauit Deum dicens uere hic homo iustus erat 23:48 et om- 23:48 And all the people that came together to that sight, beholding the things which were done, smote their breasts, and returned.

nis turba eorum qui simul aderant ad spectaculum istud et uidebant quae fiebant percutientes pectora sua reuertebantur 23:49 stabant autem omnes noti eius a longe et mulieres quae 23:49 And all his acquaintance, and the women that fol-

lowed him from Galilee, stood afar off, beholding these things.

secutae erant eum a Galilaea haec uidentes 23:50 et ecce uir

23:50 And, behold, there was a man named Joseph, a counsellor; and he was a good man, and a just:

nomine Ioseph qui erat decurio uir bonus et iustus 23:51 hic _{23:51 (The same had not consented to the counsel and}

deed of them;) he was of Arimathaea, a city of the Jews: who also himself waited for the kingdom of God.

non consenserat consilio et actibus eorum ab Arimathia ciuitate Iudaeae qui expectabat et ipse regnum Dei 23:52 hic 23:52 This man went unto Pilate, and begged the body of

Jesus.

accessit ad Pilatum et petiit corpus Iesu 23:53 et deposi- 23:53 And he took it down, and wrapped it in linen, and

laid it in a sepulchre that was hewn in stone, wherein never man before was laid.

tum inuoluit sindone et posuit eum in monumento exciso in quo nondum quisquam positus fuerat 23:54 et dies erat pa- 23:54 And that day was the preparation, and the sabbath

drew on.

rasceues et sabbatum inlucescebat 23:55 subsecutae autem 23:55 And the women also, which came with him from

Galilee, followed after, and beheld the sepulchre, and how his body was laid.

mulieres quae cum ipso uenerant de Galilaea uiderunt monu-

mentum et quemadmodum positum erat corpus eius 23:56 23:56 And they returned, and prepared spices and ointments; and rested the sabbath day according to the commandment.

et reuertentes parauerunt aromata et unguenta et sabbato quidem siluerunt secundum mandatum

24:1 una autem sabbati ualde diluculo uenerunt ad mo- 24:1 Now upon the first day of the week, very early in

the morning, they came unto the sepulchre, bringing the spices which they had prepared, and certain others with

numentum portantes quae parauerant aromata 24:2 et in- them.

24:2 And they found the stone rolled away from the sepulchre.

uenerunt lapidem reuolutum a monumento 24:3 et ingres-

24:3 And they entered in, and found not the body of the Lord Jesus.

sae non inuenerunt corpus Domini Iesu 24:4 et factum est _{24:4 And it came to pass, as they were much perplexed}

thereabout, behold, two men stood by them in shining garments:

dum mente consternatae essent de isto ecce duo uiri steterunt secus illas in ueste fulgenti 24:5 cum timerent autem 24:5 And as they were afraid, and bowed down their faces

to the earth, they said unto them, Why seek ye the living among the dead?

et declinarent uultum in terram dixerunt ad illas quid quae-

Biblia Sacra Vulgata

24:6 He is not here, but is risen: remember how he spake ritis uiuentem cum mortuis 24:6 non est hic sed surrexit re-

unto you when he was yet in Galilee,

cordamini qualiter locutus est uobis cum adhuc in Galilaea

24:7 Saying, The Son of man must be delivered into the esset 24:7 dicens quia oportet Filium hominis tradi in ma-

hands of sinful men, and be crucified, and the third day rise again.

nus hominum peccatorum et crucifigi et die tertia resurgere

24:8 And they remembered his words, 24:8 et recordatae sunt uerborum eius 24:9 et regressae a

24:9 And returned from the sepulchre, and told all these things unto the eleven, and to all the rest.

monumento nuntiauerunt haec omnia illis undecim et cete-

24:10 It was Mary Magdalene and Joanna, and Mary the ris omnibus 24:10 erat autem Maria Magdalene et Iohanna

mother of James, and other women that were with them, which told these things unto the apostles.

et Maria Iacobi et ceterae quae cum eis erant quae dicebant

24:11 And their words seemed to them as idle tales, and ad apostolos haec 24:11 et uisa sunt ante illos sicut delira-

they believed them not.

24:12 Then arose Peter, and ran unto the sepulchre; and mentum uerba ista et non credebant illis 24:12 Petrus au-

stooping down, he beheld the linen clothes laid by themselves, and departed, wondering in himself at that which

was come to pass. tem surgens cucurrit ad monumentum et procumbens uidet

linteamina sola posita et abiit secum mirans quod factum

24:13 And, behold, two of them went that same day to a fuerat 24:13 et ecce duo ex illis ibant ipsa die in castellum

village called Emmaus, which was from Jerusalem about threescore furlongs.

quod erat in spatio stadiorum sexaginta ab Hierusalem no-

24:14 And they talked together of all these things which mine Emmaus 24:14 et ipsi loquebantur ad inuicem de his had happened.

24:15 And it came to pass, that, while they communed omnibus quae acciderant 24:15 et factum est dum fabula-

together and reasoned, Jesus himself drew near, and went with them.

rentur et secum quaererent et ipse Iesus adpropinquans ibat

24:16 But their eyes were holden that they should not cum illis 24:16 oculi autem illorum tenebantur ne eum ag-

know him.

24:17 And he said unto them, What manner of communi- noscerent 24:17 et ait ad illos qui sunt hii sermones quos

cations are these that ye have one to another, as ye walk, and are sad?

confertis ad inuicem ambulantes et estis tristes 24:18 et re-

24:18 And the one of them, whose name was Cleopas, answering said unto him, Art thou only a stranger in Jerusalem, and hast not known the things which are come to pass there in these days?

spondens unus cui nomen Cleopas dixit ei tu solus peregrinus es in Hierusalem et non cognouisti quae facta sunt in

24:19 And he said unto them, What things? And they illa his diebus 24:19 quibus ille dixit quae et dixerunt de

said unto him, Concerning Jesus of Nazareth, which was a prophet mighty in deed and word before God and all the

people: Iesu Nazareno qui fuit uir propheta potens in opere et ser-

24:20 And how the chief priests and our rulers delivered mone coram Deo et omni populo 24:20 et quomodo eum him to be condemned to death, and have crucified him.

tradiderunt summi sacerdotum et principes nostri in damna-

24:21 But we trusted that it had been he which should have tionem mortis et crucifixerunt eum 24:21 nos autem spera-

redeemed Israel: and beside all this, to day is the third day since these things were done.

bamus quia ipse esset redeepturus Israhel et nunc super haec

24:22 Yea, and certain women also of our company made omnia tertia dies hodie quod haec facta sunt 24:22 sed et

us astonished, which were early at the sepulchre;

mulieres quaedam ex nostris terruerunt nos quae ante lucem

24:23 And when they found not his body, they came, say- fuerunt ad monumentum 24:23 et non inuento corpore eius

ing, that they had also seen a vision of angels, which said that he was alive.

uenerunt dicentes se etiam uisionem angelorum uidisse qui

24:24 And certain of them which were with us went to the dicunt eum uiuere 24:24 et abierunt quidam ex nostris ad

sepulchre, and found it even so as the women had said: but him they saw not.

Euangelium Secundum Lucam

monumentum et ita inuenerunt sicut mulieres dixerunt ipsum uero non uiderunt 24:25 et ipse dixit ad eos o stulti et 24:25 Then he said unto them, O fools, and slow of heart

to believe all that the prophets have spoken:

tardi corde ad credendum in omnibus quae locuti sunt prophetae 24:26 nonne haec oportuit pati Christum et ita in- 24:26 Ought not Christ to have suffered these things, and

to enter into his glory?

trare in gloriam suam 24:27 et incipiens a Mose et omnibus 24:27 And beginning at Moses and all the prophets, he

expounded unto them in all the scriptures the things concerning himself.

prophetis interpretabatur illis in omnibus scripturis quae de ipso erant 24:28 et adpropinquauerunt castello quo ibant et 24:28 And they drew nigh unto the village, whither they

went: and he made as though he would have gone further.

ipse se finxit longius ire 24:29 et coegerunt illum dicentes 24:29 But they constrained him, saying, Abide with us:

for it is toward evening, and the day is far spent. And he went in to tarry with them.

mane nobiscum quoniam aduesperascit et inclinata est iam dies et intrauit cum illis 24:30 et factum est dum recumberet 24:30 And it came to pass, as he sat at meat with them, he

took bread, and blessed it, and brake, and gave to them.

cum illis accepit panem et benedixit ac fregit et porrigebat illis 24:31 et aperti sunt oculi eorum et cognouerunt eum 24:31 And their eyes were opened, and they knew him;

and he vanished out of their sight.

et ipse euanuit ex oculis eorum 24:32 et dixerunt ad inui- 24:32 And they said one to another, Did not our heart burn

within us, while he talked with us by the way, and while he opened to us the scriptures?

cem nonne cor nostrum ardens erat in nobis dum loqueretur in uia et aperiret nobis scripturas 24:33 et surgentes eadem 24:33 And they rose up the same hour, and returned to

Jerusalem, and found the eleven gathered together, and them that were with them,

hora regressi sunt in Hierusalem et inuenerunt congregatos undecim et eos qui cum ipsis erant 24:34 dicentes quod sur- 24:34 Saying, The Lord is risen indeed, and hath appeared

to Simon.

rexit Dominus uere et apparuit Simoni 24:35 et ipsi narra- 24:35 And they told what things were done in the way, and

how he was known of them in breaking of bread.

bant quae gesta erant in uia et quomodo cognouerunt eum in fractione panis 24:36 dum haec autem loquuntur Iesus stetit 24:36 And as they thus spake, Jesus himself stood in the

midst of them, and saith unto them, Peace be unto you.

in medio eorum et dicit eis pax uobis ego sum nolite timere 24:37 conturbati uero et conterriti existimabant se spiritum 24:37 But they were terrified and affrighted, and supposed

that they had seen a spirit.

uidere 24:38 et dixit eis quid turbati estis et cogitationes as- 24:38 And he said unto them, Why are ye troubled? and

why do thoughts arise in your hearts?

cendunt in corda uestra 24:39 uidete manus meas et pedes 24:39 Behold my hands and my feet, that it is I myself:

handle me, and see; for a spirit hath not flesh and bones, as ye see me have.

quia ipse ego sum palpate et uidete quia spiritus carnem et ossa non habet sicut me uidetis habere 24:40 et cum hoc 24:40 And when he had thus spoken, he shewed them his

hands and his feet.

dixisset ostendit eis manus et pedes 24:41 adhuc autem il- 24:41 And while they yet believed not for joy, and won-

dered, he said unto them, Have ye here any meat?

lis non credentibus et mirantibus prae gaudio dixit habetis hic aliquid quod manducetur 24:42 at illi obtulerunt ei par- 24:42 And they gave him a piece of a broiled fish, and of

an honeycomb.

tem piscis assi et fauum mellis 24:43 et cum manducasset 24:43 And he took it, and did eat before them. coram eis sumens reliquias dedit eis 24:44 et dixit ad eos 24:44 And he said unto them, These are the words which

I spake unto you, while I was yet with you, that all things must be fulfilled, which were written in the law of Moses,

haec sunt uerba quae locutus sum ad uos cum adhuc essem and in the prophets, and in the psalms, concerning me. uobiscum quoniam necesse est impleri omnia quae scripta

24:45 Then opened he their understanding, that they might sunt in lege Mosi et prophetis et psalmis de me 24:45 tunc understand the scriptures,

24:46 And said unto them, Thus it is written, and thus it aperuit illis sensum ut intellegerent scripturas 24:46 et dixit

behoved Christ to suffer, and to rise from the dead the third day:

eis quoniam sic scriptum est et sic oportebat Christum pati

24:47 And that repentance and remission of sins should et resurgere a mortuis die tertia 24:47 et praedicari in no-

be preached in his name among all nations, beginning at Jerusalem.

mine eius paenitentiam et remissionem peccatorum in om-

24:48 And ye are witnesses of these things. nes gentes incipientibus ab Hierosolyma 24:48 uos autem 24:49 And, behold, I send the promise of my Father upon estis testes horum 24:49 et ego mitto promissum Patris mei

you: but tarry ye in the city of Jerusalem, until ye be endued with power from on high.

in uos uos autem sedete in ciuitate quoadusque induamini

24:50 And he led them out as far as to Bethany, and he uirtutem ex alto 24:50 eduxit autem eos foras in Bethaniam

lifted up his hands, and blessed them.

24:51 And it came to pass, while he blessed them, he was et eleuatis manibus suis benedixit eis 24:51 et factum est

parted from them, and carried up into heaven.

dum benediceret illis recessit ab eis et ferebatur in caelum

24:52 And they worshipped him, and returned to 24:52 et ipsi adorantes regressi sunt in Hierusalem cum gau-

Jerusalem with great joy:

24:53 And were continually in the temple, praising and dio magno 24:53 et erant semper in templo laudantes et be-

blessing God. Amen.

nedicentes Deum amen

Euangelium Secundum Ioannem

1:1 in principio erat Verbum et Verbum erat apud Deum 1:1 In the beginning was the Word, and the Word was with

God, and the Word was God.

et Deus erat Verbum 1:2 hoc erat in principio apud De- 1:2 The same was in the beginning with God. um 1:3 omnia per ipsum facta sunt et sine ipso factum est 1:3 All things were made by him; and without him was

not any thing made that was made.

nihil quod factum est 1:4 in ipso uita erat et uita erat lux 1:4 In him was life; and the life was the light of men. hominum 1:5 et lux in tenebris lucet et tenebrae eam non 1:5 And the light shineth in darkness; and the darkness

comprehended it not.

conprehenderunt 1:6 fuit homo missus a Deo cui nomen 1:6 There was a man sent from God, whose name was

John.

erat Iohannes 1:7 hic uenit in testimonium ut testimonium 1:7 The same came for a witness, to bear witness of the

Light, that all men through him might believe.

perhiberet de lumine ut omnes crederent per illum 1:8 non 1:8 He was not that Light, but was sent to bear witness of

that Light.

erat ille lux sed ut testimonium perhiberet de lumine 1:9 1:9 That was the true Light, which lighteth every man that

cometh into the world.

erat lux uera quae inluminat omnem hominem uenientem in

mundum 1:10 in mundo erat et mundus per ipsum factus 1:10 He was in the world, and the world was made by him,

and the world knew him not.

est et mundus eum non cognouit 1:11 in propria uenit et sui 1:11 He came unto his own, and his own received him not. eum non receperunt 1:12 quotquot autem receperunt eum 1:12 But as many as received him, to them gave he power

to become the sons of God, even to them that believe on his name:

dedit eis potestatem filios Dei fieri his qui credunt in nomine eius 1:13 qui non ex sanguinibus neque ex uoluntate carnis 1:13 Which were born, not of blood, nor of the will of the

flesh, nor of the will of man, but of God.

neque ex uoluntate uiri sed ex Deo nati sunt 1:14 et Verbum 1:14 And the Word was made flesh, and dwelt among us,

(and we beheld his glory, the glory as of the only begotten of the Father,) full of grace and truth.

caro factum est et habitauit in nobis et uidimus gloriam eius gloriam quasi unigeniti a Patre plenum gratiae et ueritatis 1:15 Iohannes testimonium perhibet de ipso et clamat dicens 1:15 John bare witness of him, and cried, saying, This was

he of whom I spake, He that cometh after me is preferred before me: for he was before me.

hic erat quem dixi uobis qui post me uenturus est ante me factus est quia prior me erat 1:16 et de plenitudine eius nos 1:16 And of his fulness have all we received, and grace for

grace.

omnes accepimus et gratiam pro gratia 1:17 quia lex per 1:17 For the law was given by Moses, but grace and truth

came by Jesus Christ.

Mosen data est gratia et ueritas per Iesum Christum facta est 1:18 Deum nemo uidit umquam unigenitus Filius qui est in 1:18 No man hath seen God at any time, the only begotten

Son, which is in the bosom of the Father, he hath declared him.

sinu Patris ipse enarrauit 1:19 et hoc est testimonium Io-

1:19 And this is the record of John, when the Jews sent priests and Levites from Jerusalem to ask him, Who art thou?

hannis quando miserunt Iudaei ab Hierosolymis sacerdotes et Leuitas ad eum ut interrogarent eum tu quis es 1:20 et 1:20 And he confessed, and denied not; but confessed, I

am not the Christ.

confessus est et non negauit et confessus est quia non sum ego Christus 1:21 et interrogauerunt eum quid ergo Helias 1:21 And they asked him, What then? Art thou Elias?

And he saith, I am not. Art thou that prophet? And he answered, No.

es tu et dicit non sum propheta es tu et respondit non 1:22

1:22 Then said they unto him, Who art thou? that we may give an answer to them that sent us. What sayest thou of thyself?

dixerunt ergo ei quis es ut responsum demus his qui miserunt nos quid dicis de te ipso 1:23 ait ego uox clamantis 1:23 He said, I am the voice of one crying in the wilder-

ness, Make straight the way of the Lord, as said the prophet Esaias.

in deserto dirigite uiam Domini sicut dixit Esaias propheta

Biblia Sacra Vulgata

1:24 And they which were sent were of the Pharisees. 1:24 et qui missi fuerant erant ex Pharisaeis 1:25 et inter-

1:25 And they asked him, and said unto him, Why baptizest thou then, if thou be not that Christ, nor Elias, neither that prophet?

rogauerunt eum et dixerunt ei quid ergo baptizas si tu non

1:26 John answered them, saying, I baptize with water: es Christus neque Helias neque propheta 1:26 respondit eis

but there standeth one among you, whom ye know not;

Iohannes dicens ego baptizo in aqua medius autem uestrum

1:27 He it is, who coming after me is preferred before me, stetit quem uos non scitis 1:27 ipse est qui post me uenturus

whose shoe’s latchet I am not worthy to unloose.

est qui ante me factus est cuius ego non sum dignus ut sol-

1:28 These things were done in Bethabara beyond Jordan, uam eius corrigiam calciamenti 1:28 haec in Bethania facta

where John was baptizing.

1:29 The next day John seeth Jesus coming unto him, and sunt trans Iordanen ubi erat Iohannes baptizans 1:29 altera

saith, Behold the Lamb of God, which taketh away the sin of the world.

die uidet Iohannes Iesum uenientem ad se et ait ecce agnus

1:30 This is he of whom I said, After me cometh a man Dei qui tollit peccatum mundi 1:30 hic est de quo dixi post

which is preferred before me: for he was before me.

1:31 And I knew him not: but that he should be made man- me uenit uir qui ante me factus est quia prior me erat 1:31

ifest to Israel, therefore am I come baptizing with water.

et ego nesciebam eum sed ut manifestaretur Israhel propte-

1:32 And John bare record, saying, I saw the Spirit de- rea ueni ego in aqua baptizans 1:32 et testimonium perhi-

scending from heaven like a dove, and it abode upon him.

buit Iohannes dicens quia uidi Spiritum descendentem quasi

1:33 And I knew him not: but he that sent me to baptize columbam de caelo et mansit super eum 1:33 et ego nescie-

with water, the same said unto me, Upon whom thou shalt see the Spirit descending, and remaining on him, the same

is he which baptizeth with the Holy Ghost. bam eum sed qui misit me baptizare in aqua ille mihi dixit

super quem uideris Spiritum descendentem et manentem su-

1:34 And I saw, and bare record that this is the Son of God. per eum hic est qui baptizat in Spiritu Sancto 1:34 et ego

1:35 Again the next day after John stood, and two of his uidi et testimonium perhibui quia hic est Filius Dei 1:35

disciples;

altera die iterum stabat Iohannes et ex discipulis eius duo

1:36 And looking upon Jesus as he walked, he saith, Be- 1:36 et respiciens Iesum ambulantem dicit ecce agnus Dei

hold the Lamb of God!

1:37 And the two disciples heard him speak, and they fol- 1:37 et audierunt eum duo discipuli loquentem et secuti sunt

lowed Jesus.

1:38 Then Jesus turned, and saw them following, and Iesum 1:38 conuersus autem Iesus et uidens eos sequen-

saith unto them, What seek ye? They said unto him, Rabbi, (which is to say, being interpreted, Master,) where

dwellest thou? tes dicit eis quid quaeritis qui dixerunt ei rabbi quod dicitur 1:39 He saith unto them, Come and see. They came and interpretatum magister ubi habitas 1:39 dicit eis uenite et

saw where he dwelt, and abode with him that day: for it was about the tenth hour.

uidete uenerunt et uiderunt ubi maneret et apud eum manse-

1:40 One of the two which heard John speak, and followed runt die illo hora autem erat quasi decima 1:40 erat autem

him, was Andrew, Simon Peter’s brother.

Andreas frater Simonis Petri unus ex duobus qui audierant

1:41 He first findeth his own brother Simon, and saith unto ab Iohanne et secuti fuerant eum 1:41 inuenit hic primum him, We have found the Messias, which is, being interpreted, the Christ.

fratrem suum Simonem et dicit ei inuenimus Messiam quod

1:42 And he brought him to Jesus. And when Jesus beheld est interpretatum Christus 1:42 et adduxit eum ad Iesum in-

him, he said, Thou art Simon the son of Jona: thou shalt be called Cephas, which is by interpretation, A stone.

tuitus autem eum Iesus dixit tu es Simon filius Iohanna tu

1:43 The day following Jesus would go forth into Galilee, uocaberis Cephas quod interpretatur Petrus 1:43 in crasti-

and findeth Philip, and saith unto him, Follow me.

Euangelium Secundum Ioannem

num uoluit exire in Galilaeam et inuenit Philippum et dicit ei Iesus sequere me 1:44 erat autem Philippus a Bethsaida ci- 1:44 Now Philip was of Bethsaida, the city of Andrew and

Peter.

uitate Andreae et Petri 1:45 inuenit Philippus Nathanahel et 1:45 Philip findeth Nathanael, and saith unto him, We have

found him, of whom Moses in the law, and the prophets, did write, Jesus of Nazareth, the son of Joseph.

dicit ei quem scripsit Moses in lege et prophetae inuenimus Iesum filium Ioseph a Nazareth 1:46 et dixit ei Nathanahel 1:46 And Nathanael said unto him, Can there any good

thing come out of Nazareth? Philip saith unto him, Come and see.

a Nazareth potest aliquid boni esse dicit ei Philippus ueni et uide 1:47 uidit Iesus Nathanahel uenientem ad se et dicit de 1:47 Jesus saw Nathanael coming to him, and saith of him,

Behold an Israelite indeed, in whom is no guile!

eo ecce uere Israhelita in quo dolus non est 1:48 dicit ei Na- 1:48 Nathanael saith unto him, Whence knowest thou

me? Jesus answered and said unto him, Before that Philip called thee, when thou wast under the fig tree, I saw thee.

thanahel unde me nosti respondit Iesus et dixit ei priusquam te Philippus uocaret cum esses sub ficu uidi te 1:49 respon- 1:49 Nathanael answered and saith unto him, Rabbi, thou

art the Son of God; thou art the King of Israel.

dit ei Nathanahel et ait rabbi tu es Filius Dei tu es rex Israhel 1:50 respondit Iesus et dixit ei quia dixi tibi uidi te sub ficu 1:50 Jesus answered and said unto him, Because I said

unto thee, I saw thee under the fig tree, believest thou? thou shalt see greater things than these.

credis maius his uidebis 1:51 et dicit ei amen amen dico

1:51 And he saith unto him, Verily, verily, I say unto you, Hereafter ye shall see heaven open, and the angels of God ascending and descending upon the Son of man.

uobis uidebitis caelum apertum et angelos Dei ascendentes et descendentes supra Filium hominis

2:1 et die tertio nuptiae factae sunt in Cana Galilaeae et 2:1 And the third day there was a marriage in Cana of

Galilee; and the mother of Jesus was there:

erat mater Iesu ibi 2:2 uocatus est autem ibi et Iesus et dis- 2:2 And both Jesus was called, and his disciples, to the

marriage.

cipuli eius ad nuptias 2:3 et deficiente uino dicit mater Iesu 2:3 And when they wanted wine, the mother of Jesus saith

unto him, They have no wine.

ad eum uinum non habent 2:4 et dicit ei Iesus quid mihi et 2:4 Jesus saith unto her, Woman, what have I to do with

thee? mine hour is not yet come.

tibi est mulier nondum uenit hora mea 2:5 dicit mater eius 2:5 His mother saith unto the servants, Whatsoever he

saith unto you, do it.

ministris quodcumque dixerit uobis facite 2:6 erant autem 2:6 And there were set there six waterpots of stone, after

the manner of the purifying of the Jews, containing two or three firkins apiece.

ibi lapideae hydriae sex positae secundum purificationem Iudaeorum capientes singulae metretas binas uel ternas 2:7 2:7 Jesus saith unto them, Fill the waterpots with water.

And they filled them up to the brim.

dicit eis Iesus implete hydrias aqua et impleuerunt eas usque ad summum 2:8 et dicit eis Iesus haurite nunc et ferte 2:8 And he saith unto them, Draw out now, and bear unto

the governor of the feast. And they bare it.

architriclino et tulerunt 2:9 ut autem gustauit architriclinus 2:9 When the ruler of the feast had tasted the water that

was made wine, and knew not whence it was: (but the servants which drew the water knew;) the governor of the

aquam uinum factam et non sciebat unde esset ministri au- feast called the bridegroom, tem sciebant qui haurierant aquam uocat sponsum architri-

clinus 2:10 et dicit ei omnis homo primum bonum uinum 2:10 And saith unto him, Every man at the beginning doth set forth good wine; and when men have well drunk, then that which is worse: but thou hast kept the good wine until

ponit et cum inebriati fuerint tunc id quod deterius est tu now. seruasti bonum uinum usque adhuc 2:11 hoc fecit initium 2:11 This beginning of miracles did Jesus in Cana of

Galilee, and manifested forth his glory; and his disciples believed on him.

signorum Iesus in Cana Galilaeae et manifestauit gloriam suam et crediderunt in eum discipuli eius 2:12 post hoc de- 2:12 After this he went down to Capernaum, he, and his

mother, and his brethren, and his disciples: and they continued there not many days.

Biblia Sacra Vulgata

scendit Capharnaum ipse et mater eius et fratres eius et dis-

2:13 And the Jews’ passover was at hand, and Jesus went cipuli eius et ibi manserunt non multis diebus 2:13 et prope

up to Jerusalem.

2:14 And found in the temple those that sold oxen and erat pascha Iudaeorum et ascendit Hierosolyma Iesus 2:14

sheep and doves, and the changers of money sitting:

et inuenit in templo uendentes boues et oues et columbas

2:15 And when he had made a scourge of small cords, he et nummularios sedentes 2:15 et cum fecisset quasi flagel-

drove them all out of the temple, and the sheep, and the oxen; and poured out the changers’ money, and overthrew

the tables; lum de funiculis omnes eiecit de templo oues quoque et bo-2:16 And said unto them that sold doves, Take these things ues et nummulariorum effudit aes et mensas subuertit 2:16

hence; make not my Father’s house an house of merchandise.

et his qui columbas uendebant dixit auferte ista hinc nolite

2:17 And his disciples remembered that it was written, The facere domum Patris mei domum negotiationis 2:17 recorzeal of thine house hath eaten me up.

dati uero sunt discipuli eius quia scriptum est zelus domus

2:18 Then answered the Jews and said unto him, What tuae comedit me 2:18 responderunt ergo Iudaei et dixerunt

sign shewest thou unto us, seeing that thou doest these things?

ei quod signum ostendis nobis quia haec facis 2:19 respon-

2:19 Jesus answered and said unto them, Destroy this temple, and in three days I will raise it up.

dit Iesus et dixit eis soluite templum hoc et in tribus diebus

2:20 Then said the Jews, Forty and six years was this tem- excitabo illud 2:20 dixerunt ergo Iudaei quadraginta et sex

ple in building, and wilt thou rear it up in three days?

annis aedificatum est templum hoc et tu tribus diebus exci-

2:21 But he spake of the temple of his body. tabis illud 2:21 ille autem dicebat de templo corporis sui 2:22 When therefore he was risen from the dead, his disci- 2:22 cum ergo resurrexisset a mortuis recordati sunt disci-

ples remembered that he had said this unto them; and they believed the scripture, and the word which Jesus had said.

puli eius quia hoc dicebat et crediderunt scripturae et ser-

2:23 Now when he was in Jerusalem at the passover, in the moni quem dixit Iesus 2:23 cum autem esset Hierosolymis

feast day, many believed in his name, when they saw the miracles which he did.

in pascha in die festo multi crediderunt in nomine eius ui-

2:24 But Jesus did not commit himself unto them, because dentes signa eius quae faciebat 2:24 ipse autem Iesus non

he knew all men, 2:25 And needed not that any should testify of man: for he credebat semet ipsum eis eo quod ipse nosset omnes 2:25 knew what was in man.

et quia opus ei non erat ut quis testimonium perhiberet de homine ipse enim sciebat quid esset in homine

3:1 There was a man of the Pharisees, named Nicodemus, 3:1 erat autem homo ex Pharisaeis Nicodemus nomine a ruler of the Jews:

3:2 The same came to Jesus by night, and said unto him, princeps Iudaeorum 3:2 hic uenit ad eum nocte et dixit ei

Rabbi, we know that thou art a teacher come from God: for no man can do these miracles that thou doest, except

God be with him. rabbi scimus quia a Deo uenisti magister nemo enim po-

test haec signa facere quae tu facis nisi fuerit Deus cum eo

3:3 Jesus answered and said unto him, Verily, verily, I say 3:3 respondit Iesus et dixit ei amen amen dico tibi nisi quis

unto thee, Except a man be born again, he cannot see the kingdom of God.

natus fuerit denuo non potest uidere regnum Dei 3:4 dicit

3:4 Nicodemus saith unto him, How can a man be born when he is old? can he enter the second time into his mother’s womb, and be born?

ad eum Nicodemus quomodo potest homo nasci cum senex sit numquid potest in uentrem matris suae iterato introire et

3:5 Jesus answered, Verily, verily, I say unto thee, Except nasci 3:5 respondit Iesus amen amen dico tibi nisi quis rea man be born of water and of the Spirit, he cannot enter into the kingdom of God.

Euangelium Secundum Ioannem

natus fuerit ex aqua et Spiritu non potest introire in regnum Dei 3:6 quod natum est ex carne caro est et quod natum est 3:6 That which is born of the flesh is flesh; and that which

is born of the Spirit is spirit.

ex Spiritu spiritus est 3:7 non mireris quia dixi tibi oportet 3:7 Marvel not that I said unto thee, Ye must be born again. uos nasci denuo 3:8 Spiritus ubi uult spirat et uocem eius 3:8 The wind bloweth where it listeth, and thou hearest

the sound thereof, but canst not tell whence it cometh, and whither it goeth: so is every one that is born of the Spirit.

audis sed non scis unde ueniat et quo uadat sic est omnis qui natus est ex Spiritu 3:9 respondit Nicodemus et dixit 3:9 Nicodemus answered and said unto him, How can

these things be?

ei quomodo possunt haec fieri 3:10 respondit Iesus et dixit 3:10 Jesus answered and said unto him, Art thou a master

of Israel, and knowest not these things?

ei tu es magister Israhel et haec ignoras 3:11 amen amen 3:11 Verily, verily, I say unto thee, We speak that we do

know, and testify that we have seen; and ye receive not our witness.

dico tibi quia quod scimus loquimur et quod uidimus testamur et testimonium nostrum non accipitis 3:12 si terrena 3:12 If I have told you earthly things, and ye believe not,

how shall ye believe, if I tell you of heavenly things?

dixi uobis et non creditis quomodo si dixero uobis caelestia credetis 3:13 et nemo ascendit in caelum nisi qui descendit 3:13 And no man hath ascended up to heaven, but he that

came down from heaven, even the Son of man which is in heaven.

de caelo Filius hominis qui est in caelo 3:14 et sicut Mo-

3:14 And as Moses lifted up the serpent in the wilderness, even so must the Son of man be lifted up:

ses exaltauit serpentem in deserto ita exaltari oportet Filium hominis 3:15 ut omnis qui credit in ipso non pereat sed ha- 3:15 That whosoever believeth in him should not perish,

but have eternal life.

beat uitam aeternam 3:16 sic enim dilexit Deus mundum ut 3:16 For God so loved the world, that he gave his only

begotten Son, that whosoever believeth in him should not perish, but have everlasting life.

Filium suum unigenitum daret ut omnis qui credit in eum non pereat sed habeat uitam aeternam 3:17 non enim mi- 3:17 For God sent not his Son into the world to condemn

the world; but that the world through him might be saved.

sit Deus Filium suum in mundum ut iudicet mundum sed ut saluetur mundus per ipsum 3:18 qui credit in eum non iudi- 3:18 He that believeth on him is not condemned: but he

that believeth not is condemned already, because he hath not believed in the name of the only begotten Son of God.

catur qui autem non credit iam iudicatus est quia non credidit in nomine unigeniti Filii Dei 3:19 hoc est autem iudicium 3:19 And this is the condemnation, that light is come into

the world, and men loved darkness rather than light, because their deeds were evil.

quia lux uenit in mundum et dilexerunt homines magis tene-

bras quam lucem erant enim eorum mala opera 3:20 omnis 3:20 For every one that doeth evil hateth the light, neither

cometh to the light, lest his deeds should be reproved.

enim qui mala agit odit lucem et non uenit ad lucem ut non arguantur opera eius 3:21 qui autem facit ueritatem uenit 3:21 But he that doeth truth cometh to the light, that his

deeds may be made manifest, that they are wrought in God.

ad lucem ut manifestentur eius opera quia in Deo sunt facta 3:22 post haec uenit Iesus et discipuli eius in iudaeam ter- 3:22 After these things came Jesus and his disciples into

the land of Judaea; and there he tarried with them, and baptized.

ram et illic demorabatur cum eis et baptizabat 3:23 erat au-

3:23 And John also was baptizing in Aenon near to Salim, because there was much water there: and they came, and were baptized.

tem et Iohannes baptizans in Aenon iuxta Salim quia aquae multae erant illic et adueniebant et baptizabantur 3:24 non- 3:24 For John was not yet cast into prison. dum enim missus fuerat in carcerem Iohannes 3:25 facta 3:25 Then there arose a question between some of John’s

disciples and the Jews about purifying.

est ergo quaestio ex discipulis Iohannis cum Iudaeis de puri-

Biblia Sacra Vulgata

3:26 And they came unto John, and said unto him, Rabbi, ficatione 3:26 et uenerunt ad Iohannem et dixerunt ei rabbi

he that was with thee beyond Jordan, to whom thou barest witness, behold, the same baptizeth, and all men come to

him. qui erat tecum trans Iordanen cui tu testimonium perhibuisti 3:27 John answered and said, A man can receive nothing, ecce hic baptizat et omnes ueniunt ad eum 3:27 respondit

except it be given him from heaven.

Iohannes et dixit non potest homo accipere quicquam nisi

3:28 Ye yourselves bear me witness, that I said, I am not fuerit ei datum de caelo 3:28 ipsi uos mihi testimonium per-

the Christ, but that I am sent before him.

hibetis quod dixerim ego non sum Christus sed quia missus

3:29 He that hath the bride is the bridegroom: but the sum ante illum 3:29 qui habet sponsam sponsus est ami-

friend of the bridegroom, which standeth and heareth him, rejoiceth greatly because of the bridegroom’s voice: this

my joy therefore is fulfilled. cus autem sponsi qui stat et audit eum gaudio gaudet propter 3:30 He must increase, but I must decrease. uocem sponsi hoc ergo gaudium meum impletum est 3:30 3:31 He that cometh from above is above all: he that is illum oportet crescere me autem minui 3:31 qui desursum

of the earth is earthly, and speaketh of the earth: he that cometh from heaven is above all.

uenit supra omnes est qui est de terra de terra est et de terra

3:32 And what he hath seen and heard, that he testifieth; loquitur qui de caelo uenit supra omnes est 3:32 et quod and no man receiveth his testimony.

uidit et audiuit hoc testatur et testimonium eius nemo ac-

3:33 He that hath received his testimony hath set to his cipit 3:33 qui accipit eius testimonium signauit quia Deus

seal that God is true.

3:34 For he whom God hath sent speaketh the words of uerax est 3:34 quem enim misit Deus uerba Dei loquitur

God: for God giveth not the Spirit by measure unto him. 3:35 The Father loveth the Son, and hath given all things non enim ad mensuram dat Deus Spiritum 3:35 Pater diliinto his hand. 3:36 He that believeth on the Son hath everlasting life: git Filium et omnia dedit in manu eius 3:36 qui credit in and he that believeth not the Son shall not see life; but the wrath of God abideth on him.

Filium habet uitam aeternam qui autem incredulus est Filio non uidebit uitam sed ira Dei manet super eum

4:1 When therefore the LORD knew how the Pharisees 4:1 ut ergo cognouit Iesus quia audierunt Pharisaei quia

had heard that Jesus made and baptized more disciples than John,

Iesus plures discipulos facit et baptizat quam Iohannes 4:2

4:2 (Though Jesus himself baptized not, but his disciples,)

4:3 He left Judaea, and departed again into Galilee. quamquam Iesus non baptizaret sed discipuli eius 4:3 re-4:4 And he must needs go through Samaria. liquit Iudaeam et abiit iterum in Galilaeam 4:4 oportebat 4:5 Then cometh he to a city of Samaria, which is called autem eum transire per Samariam 4:5 uenit ergo in ciuita-

Sychar, near to the parcel of ground that Jacob gave to his son Joseph.

tem Samariae quae dicitur Sychar iuxta praedium quod dedit

4:6 Now Jacob’s well was there. Jesus therefore, being Iacob Ioseph filio suo 4:6 erat autem ibi fons Iacob Iesus

wearied with his journey, sat thus on the well: and it was about the sixth hour.

ergo fatigatus ex itinere sedebat sic super fontem hora erat

4:7 There cometh a woman of Samaria to draw water: Je- quasi sexta 4:7 uenit mulier de Samaria haurire aquam di-

sus saith unto her, Give me to drink.

4:8 (For his disciples were gone away unto the city to buy cit ei Iesus da mihi bibere 4:8 discipuli enim eius abierant

meat.)

4:9 Then saith the woman of Samaria unto him, How is in ciuitatem ut cibos emerent 4:9 dicit ergo ei mulier illa

it that thou, being a Jew, askest drink of me, which am a woman of Samaria? for the Jews have no dealings with

the Samaritans. samaritana quomodo tu Iudaeus cum sis bibere a me pos-

cis quae sum mulier samaritana non enim coutuntur Iudaei

4:10 Jesus answered and said unto her, If thou knewest Samaritanis 4:10 respondit Iesus et dixit ei si scires donum

the gift of God, and who it is that saith to thee, Give me to drink; thou wouldest have asked of him, and he would have given thee living water.

Euangelium Secundum Ioannem

Dei et quis est qui dicit tibi da mihi bibere tu forsitan petisses ab eo et dedisset tibi aquam uiuam 4:11 dicit ei mulier Do- 4:11 The woman saith unto him, Sir, thou hast nothing to

draw with, and the well is deep: from whence then hast thou that living water?

mine neque in quo haurias habes et puteus altus est unde ergo

habes aquam uiuam 4:12 numquid tu maior es patre nostro 4:12 Art thou greater than our father Jacob, which gave us the well, and drank thereof himself, and his children, and his cattle?

Iacob qui dedit nobis puteum et ipse ex eo bibit et filii eius et pecora eius 4:13 respondit Iesus et dixit ei omnis qui bibit 4:13 Jesus answered and said unto her, Whosoever drin-

keth of this water shall thirst again:

ex aqua hac sitiet iterum qui autem biberit ex aqua quam ego

dabo ei non sitiet in aeternum 4:14 sed aqua quam dabo ei 4:14 But whosoever drinketh of the water that I shall give him shall never thirst; but the water that I shall give him shall be in him a well of water springing up into everlast-

fiet in eo fons aquae salientis in uitam aeternam 4:15 dicit ing life.

4:15 The woman saith unto him, Sir, give me this water, that I thirst not, neither come hither to draw.

ad eum mulier Domine da mihi hanc aquam ut non sitiam neque ueniam huc haurire 4:16 dicit ei Iesus uade uoca uirum 4:16 Jesus saith unto her, Go, call thy husband, and come

hither.

tuum et ueni huc 4:17 respondit mulier et dixit non habeo 4:17 The woman answered and said, I have no husband.

Jesus said unto her, Thou hast well said, I have no husband:

uirum dicit ei Iesus bene dixisti quia non habeo uirum 4:18

4:18 For thou hast had five husbands; and he whom thou now hast is not thy husband: in that saidst thou truly.

quinque enim uiros habuisti et nunc quem habes non est tuus uir hoc uere dixisti 4:19 dicit ei mulier Domine uideo quia 4:19 The woman saith unto him, Sir, I perceive that thou

art a prophet.

propheta es tu 4:20 patres nostri in monte hoc adorauerunt 4:20 Our fathers worshipped in this mountain; and ye say,

that in Jerusalem is the place where men ought to worship.

et uos dicitis quia Hierosolymis est locus ubi adorare oportet 4:21 dicit ei Iesus mulier crede mihi quia ueniet hora quando 4:21 Jesus saith unto her, Woman, believe me, the hour

cometh, when ye shall neither in this mountain, nor yet at Jerusalem, worship the Father.

neque in monte hoc neque in Hierosolymis adorabitis Patrem 4:22 uos adoratis quod nescitis nos adoramus quod sci- 4:22 Ye worship ye know not what: we know what we

worship: for salvation is of the Jews.

mus quia salus ex Iudaeis est 4:23 sed uenit hora et nunc est 4:23 But the hour cometh, and now is, when the true wor-

shippers shall worship the Father in spirit and in truth: for the Father seeketh such to worship him.

quando ueri adoratores adorabunt Patrem in spiritu et ueritate nam et Pater tales quaerit qui adorent eum 4:24 spiritus 4:24 God is a Spirit: and they that worship him must wor-

ship him in spirit and in truth.

est Deus et eos qui adorant eum in spiritu et ueritate oportet adorare 4:25 dicit ei mulier scio quia Messias uenit qui di- 4:25 The woman saith unto him, I know that Messias

cometh, which is called Christ: when he is come, he will tell us all things.

citur Christus cum ergo uenerit ille nobis adnuntiabit omnia 4:26 dicit ei Iesus ego sum qui loquor tecum 4:27 et conti- 4:26 Jesus saith unto her, I that speak unto thee am he.

4:27 And upon this came his disciples, and marvelled that he talked with the woman: yet no man said, What seekest thou? or, Why talkest thou with her?

nuo uenerunt discipuli eius et mirabantur quia cum muliere loquebatur nemo tamen dixit quid quaeris aut quid loqueris cum ea 4:28 reliquit ergo hydriam suam mulier et abiit 4:28 The woman then left her waterpot, and went her way

into the city, and saith to the men,

in ciuitatem et dicit illis hominibus 4:29 uenite uidete ho- 4:29 Come, see a man, which told me all things that ever I

did: is not this the Christ?

minem qui dixit mihi omnia quaecumque feci numquid ipse est Christus 4:30 exierunt de ciuitate et ueniebant ad eum 4:30 Then they went out of the city, and came unto him.

Biblia Sacra Vulgata

4:31 In the mean while his disciples prayed him, saying, 4:31 interea rogabant eum discipuli dicentes rabbi manduca

Master, eat.

4:32 But he said unto them, I have meat to eat that ye know 4:32 ille autem dixit eis ego cibum habeo manducare quem

not of. 4:33 Therefore said the disciples one to another, Hath any uos nescitis 4:33 dicebant ergo discipuli ad inuicem numman brought him ought to eat?

4:34 Jesus saith unto them, My meat is to do the will of quid aliquis adtulit ei manducare 4:34 dicit eis Iesus meus

him that sent me, and to finish his work.

cibus est ut faciam uoluntatem eius qui misit me ut perfi-

4:35 Say not ye, There are yet four months, and then ciam opus eius 4:35 nonne uos dicitis quod adhuc quattuor

cometh harvest? behold, I say unto you, Lift up your eyes, and look on the fields; for they are white already to har-

vest. menses sunt et messis uenit ecce dico uobis leuate oculos ue-4:36 And he that reapeth receiveth wages, and gathereth stros et uidete regiones quia albae sunt iam ad messem 4:36

fruit unto life eternal: that both he that soweth and he that reapeth may rejoice together.

et qui metit mercedem accipit et congregat fructum in uitam

4:37 And herein is that saying true, One soweth, and an- aeternam ut et qui seminat simul gaudeat et qui metit 4:37

other reapeth.

in hoc enim est uerbum uerum quia alius est qui seminat et

4:38 I sent you to reap that whereon ye bestowed no alius est qui metit 4:38 ego misi uos metere quod uos non

labour: other men laboured, and ye are entered into their labours.

laborastis alii laborauerunt et uos in laborem eorum introistis

4:39 And many of the Samaritans of that city believed on 4:39 ex ciuitate autem illa multi crediderunt in eum Samari-

him for the saying of the woman, which testified, He told me all that ever I did.

tanorum propter uerbum mulieris testimonium perhibentis

4:40 So when the Samaritans were come unto him, they quia dixit mihi omnia quaecumque feci 4:40 cum uenissent

besought him that he would tarry with them: and he abode there two days.

ergo ad illum Samaritani rogauerunt eum ut ibi maneret et

4:41 And many more believed because of his own word; mansit ibi duos dies 4:41 et multo plures crediderunt pro-4:42 And said unto the woman, Now we believe, not be- pter sermonem eius 4:42 et mulieri dicebant quia iam non

cause of thy saying: for we have heard him ourselves, and know that this is indeed the Christ, the Saviour of the

world. propter tuam loquellam credimus ipsi enim audiuimus et sci-

4:43 Now after two days he departed thence, and went into mus quia hic est uere saluator mundi 4:43 post duos autem Galilee.

4:44 For Jesus himself testified, that a prophet hath no dies exiit inde et abiit in Galilaeam 4:44 ipse enim Iesus

honour in his own country.

testimonium perhibuit quia propheta in sua patria honorem

4:45 Then when he was come into Galilee, the Galilaeans non habet 4:45 cum ergo uenisset in Galilaeam exceperunt

received him, having seen all the things that he did at Jerusalem at the feast: for they also went unto the feast.

eum Galilaei cum omnia uidissent quae fecerat Hierosoly-

4:46 So Jesus came again into Cana of Galilee, where he mis in die festo et ipsi enim uenerant in diem festum 4:46

made the water wine. And there was a certain nobleman, whose son was sick at Capernaum.

uenit ergo iterum in Cana Galilaeae ubi fecit aquam uinum et erat quidam regulus cuius filius infirmabatur Capharnaum

4:47 When he heard that Jesus was come out of Judaea 4:47 hic cum audisset quia Iesus adueniret a Iudaea in Gali-

into Galilee, he went unto him, and besought him that he would come down, and heal his son: for he was at the

point of death. laeam abiit ad eum et rogabat eum ut descenderet et sanaret 4:48 Then said Jesus unto him, Except ye see signs and filium eius incipiebat enim mori 4:48 dixit ergo Iesus ad

wonders, ye will not believe.

4:49 The nobleman saith unto him, Sir, come down ere my eum nisi signa et prodiiia uideritis non creditis 4:49 dicit

child die.

ad eum regulus Domine descende priusquam moriatur filius

Euangelium Secundum Ioannem

meus 4:50 dicit ei Iesus uade filius tuus uiuit credidit homo 4:50 Jesus saith unto him, Go thy way; thy son liveth. And

the man believed the word that Jesus had spoken unto him, and he went his way.

sermoni quem dixit ei Iesus et ibat 4:51 iam autem eo de-

4:51 And as he was now going down, his servants met him, and told him, saying, Thy son liveth.

scendente serui occurrerunt ei et nuntiauerunt dicentes quia filius eius uiueret 4:52 interrogabat ergo horam ab eis in 4:52 Then enquired he of them the hour when he began to

amend. And they said unto him, Yesterday at the seventh hour the fever left him.

qua melius habuerit et dixerunt ei quia heri hora septima reliquit eum febris 4:53 cognouit ergo pater quia illa hora erat 4:53 So the father knew that it was at the same hour, in the

which Jesus said unto him, Thy son liveth: and himself believed, and his whole house.

in qua dixit ei Iesus filius tuus uiuit et credidit ipse et domus eius tota 4:54 hoc iterum secundum signum fecit Iesus cum 4:54 This is again the second miracle that Jesus did, when

he was come out of Judaea into Galilee.

uenisset a Iudaea in Galilaeam

5:1 post haec erat dies festus Iudaeorum et ascendit Ie- 5:1 After this there was a feast of the Jews; and Jesus went

up to Jerusalem.

sus Hierosolymis 5:2 est autem Hierosolymis super Pro- 5:2 Now there is at Jerusalem by the sheep market a pool,

which is called in the Hebrew tongue Bethesda, having five porches.

batica piscina quae cognominatur hebraice Bethsaida quinque porticus habens 5:3 in his iacebat multitudo magna 5:3 In these lay a great multitude of impotent folk, of

blind, halt, withered, waiting for the moving of the water.

languentium caecorum claudorum aridorum expectantium aquae motum 5:4 5:5 erat autem quidam homo ibi triginta 5:4 For an angel went down at a certain season into the

pool, and troubled the water: whosoever then first after the troubling of the water stepped in was made whole of whatsoever disease he had. 5:5 And a certain man was there, which had an infirmity thirty and eight years.

et octo annos habens in infirmitate sua 5:6 hunc cum uidis-

5:6 When Jesus saw him lie, and knew that he had been now a long time in that case, he saith unto him, Wilt thou be made whole?

set Iesus iacentem et cognouisset quia multum iam tempus habet dicit ei uis sanus fieri 5:7 respondit ei languidus Do-

5:7 The impotent man answered him, Sir, I have no man, when the water is troubled, to put me into the pool: but while I am coming, another steppeth down before me.

mine hominem non habeo ut cum turbata fuerit aqua mittat me in piscinam dum uenio enim ego alius ante me descendit 5:8 dicit ei Iesus surge tolle grabattum tuum et ambula 5:9 5:8 Jesus saith unto him, Rise, take up thy bed, and walk.

5:9 And immediately the man was made whole, and took up his bed, and walked: and on the same day was the sabbath.

et statim sanus factus est homo et sustulit grabattum suum et ambulabat erat autem sabbatum in illo die 5:10 dicebant 5:10 The Jews therefore said unto him that was cured, It is

the sabbath day: it is not lawful for thee to carry thy bed.

Iudaei illi qui sanatus fuerat sabbatum est non licet tibi tollere grabattum tuum 5:11 respondit eis qui me fecit sanum 5:11 He answered them, He that made me whole, the same

said unto me, Take up thy bed, and walk.

ille mihi dixit tolle grabattum tuum et ambula 5:12 inter- 5:12 Then asked they him, What man is that which said

unto thee, Take up thy bed, and walk?

rogauerunt ergo eum quis est ille homo qui dixit tibi tolle grabattum tuum et ambula 5:13 is autem qui sanus fuerat 5:13 And he that was healed wist not who it was: for Je-

sus had conveyed himself away, a multitude being in that place.

effectus nesciebat quis esset Iesus enim declinauit turba constituta in loco 5:14 postea inuenit eum Iesus in templo et 5:14 Afterward Jesus findeth him in the temple, and said

unto him, Behold, thou art made whole: sin no more, lest a worse thing come unto thee.

dixit illi ecce sanus factus es iam noli peccare ne deterius tibi aliquid contingat 5:15 abiit ille homo et nuntiauit Iu- 5:15 The man departed, and told the Jews that it was Jesus,

which had made him whole.

daeis quia Iesus esset qui fecit eum sanum 5:16 propterea 5:16 And therefore did the Jews persecute Jesus, and

sought to slay him, because he had done these things on the sabbath day.

Biblia Sacra Vulgata

persequebantur Iudaei Iesum quia haec faciebat in sabbato

5:17 But Jesus answered them, My Father worketh hith- 5:17 Iesus autem respondit eis Pater meus usque modo ope-

erto, and I work.

5:18 Therefore the Jews sought the more to kill him, be- ratur et ego operor 5:18 propterea ergo magis quaerebant

cause he not only had broken the sabbath, but said also that God was his Father, making himself equal with God.

eum Iudaei interficere quia non solum soluebat sabbatum sed et Patrem suum dicebat Deum aequalem se faciens Deo

5:19 Then answered Jesus and said unto them, Verily, ver- respondit itaque Iesus et dixit eis 5:19 amen amen dico uo-

ily, I say unto you, The Son can do nothing of himself, but what he seeth the Father do: for what things soever he

doeth, these also doeth the Son likewise. bis non potest Filius a se facere quicquam nisi quod uiderit

Patrem facientem quaecumque enim ille fecerit haec et Fi-

5:20 For the Father loveth the Son, and sheweth him all lius similiter facit 5:20 Pater enim diligit Filium et omnia

things that himself doeth: and he will shew him greater works than these, that ye may marvel.

demonstrat ei quae ipse facit et maiora his demonstrabit ei

5:21 For as the Father raiseth up the dead, and quickeneth opera ut uos miremini 5:21 sicut enim Pater suscitat morthem; even so the Son quickeneth whom he will.

5:22 For the Father judgeth no man, but hath committed tuos et uiuificat sic et Filius quos uult uiuificat 5:22 neque

all judgment unto the Son:

enim Pater iudicat quemquam sed iudicium omne dedit Filio

5:23 That all men should honour the Son, even as they 5:23 ut omnes honorificent Filium sicut honorificant Patrem

honour the Father. He that honoureth not the Son honoureth not the Father which hath sent him.

qui non honorificat Filium non honorificat Patrem qui misit

5:24 Verily, verily, I say unto you, He that heareth my illum 5:24 amen amen dico uobis quia qui uerbum meum word, and believeth on him that sent me, hath everlasting life, and shall not come into condemnation; but is passed

from death unto life. audit et credit ei qui misit me habet uitam aeternam et in 5:25 Verily, verily, I say unto you, The hour is coming, iudicium non uenit sed transit a morte in uitam 5:25 amen

and now is, when the dead shall hear the voice of the Son of God: and they that hear shall live.

amen dico uobis quia uenit hora et nunc est quando mortui

5:26 For as the Father hath life in himself; so hath he given audient uocem Filii Dei et qui audierint uiuent 5:26 sicut to the Son to have life in himself;

enim Pater habet uitam in semet ipso sic dedit et Filio uitam

5:27 And hath given him authority to execute judgment habere in semet ipso 5:27 et potestatem dedit ei et iudicium

also, because he is the Son of man.

5:28 Marvel not at this: for the hour is coming, in the facere quia Filius hominis est 5:28 nolite mirari hoc quia

which all that are in the graves shall hear his voice,

uenit hora in qua omnes qui in monumentis sunt audient uo-

5:29 And shall come forth; they that have done good, unto cem eius 5:29 et procedent qui bona fecerunt in resurrec-

the resurrection of life; and they that have done evil, unto the resurrection of damnation.

tionem uitae qui uero mala egerunt in resurrectionem iudi-

5:30 I can of mine own self do nothing: as I hear, I judge: cii 5:30 non possum ego a me ipso facere quicquam sicut and my judgment is just; because I seek not mine own will, but the will of the Father which hath sent me.

audio iudico et iudicium meum iustum est quia non quaero

5:31 If I bear witness of myself, my witness is not true. uoluntatem meam sed uoluntatem eius qui misit me 5:31

si ego testimonium perhibeo de me testimonium meum non

5:32 There is another that beareth witness of me; and I est uerum 5:32 alius est qui testimonium perhibet de me et

know that the witness which he witnesseth of me is true. 5:33 Ye sent unto John, and he bare witness unto the truth. scio quia uerum est testimonium quod perhibet de me 5:33

uos misistis ad Iohannem et testimonium perhibuit ueritati

Euangelium Secundum Ioannem

5:34 ego autem non ab homine testimonium accipio sed haec 5:34 But I receive not testimony from man: but these

things I say, that ye might be saved.

dico ut uos salui sitis 5:35 ille erat lucerna ardens et lucens 5:35 He was a burning and a shining light: and ye were

willing for a season to rejoice in his light.

uos autem uoluistis exultare ad horam in luce eius 5:36 ego 5:36 But I have greater witness than that of John: for the

works which the Father hath given me to finish, the same works that I do, bear witness of me, that the Father hath

autem habeo testimonium maius Iohanne opera enim quae sent me. dedit mihi Pater ut perficiam ea ipsa opera quae ego facio testimonium perhibent de me quia Pater me misit 5:37 et qui 5:37 And the Father himself, which hath sent me, hath

borne witness of me. Ye have neither heard his voice at any time, nor seen his shape.

misit me Pater ipse testimonium perhibuit de me neque uocem eius umquam audistis neque speciem eius uidistis 5:38 5:38 And ye have not his word abiding in you: for whom

he hath sent, him ye believe not.

et uerbum eius non habetis in uobis manens quia quem misit ille huic uos non creditis 5:39 scrutamini scripturas quia 5:39 Search the scriptures; for in them ye think ye have

eternal life: and they are they which testify of me.

uos putatis in ipsis uitam aeternam habere et illae sunt quae testimonium perhibent de me 5:40 et non uultis uenire ad 5:40 And ye will not come to me, that ye might have life. me ut uitam habeatis 5:41 claritatem ab hominibus non ac- 5:41 I receive not honour from men. cipio 5:42 sed cognoui uos quia dilectionem Dei non ha- 5:42 But I know you, that ye have not the love of God in

you.

betis in uobis 5:43 ego ueni in nomine Patris mei et non 5:43 I am come in my Father’s name, and ye receive me

not: if another shall come in his own name, him ye will receive.

accipitis me si alius uenerit in nomine suo illum accipietis 5:44 quomodo potestis uos credere qui gloriam ab inuicem 5:44 How can ye believe, which receive honour one of

another, and seek not the honour that cometh from God only?

accipitis et gloriam quae a solo est Deo non quaeritis 5:45

5:45 Do not think that I will accuse you to the Father: there is one that accuseth you, even Moses, in whom ye trust.

nolite putare quia ego accusaturus sim uos apud Patrem est qui accuset uos Moses in quo uos speratis 5:46 si enim cre- 5:46 For had ye believed Moses, ye would have believed

me; for he wrote of me.

deretis Mosi crederetis forsitan et mihi de me enim ille scripsit 5:47 si autem illius litteris non creditis quomodo meis 5:47 But if ye believe not his writings, how shall ye believe

my words?

uerbis credetis

6:1 post haec abiit Iesus trans mare Galilaeae quod est Ti- 6:1 After these things Jesus went over the sea of Galilee,

which is the sea of Tiberias.

beriadis 6:2 et sequebatur eum multitudo magna quia ui- 6:2 And a great multitude followed him, because they saw

his miracles which he did on them that were diseased.

debant signa quae faciebat super his qui infirmabantur 6:3 6:3 And Jesus went up into a mountain, and there he sat

with his disciples.

subiit ergo in montem Iesus et ibi sedebat cum discipulis suis 6:4 erat autem proximum pascha dies festus Iudaeorum 6:5 6:4 And the passover, a feast of the Jews, was nigh.

6:5 When Jesus then lifted up his eyes, and saw a great company come unto him, he saith unto Philip, Whence shall we buy bread, that these may eat?

cum subleuasset ergo oculos Iesus et uidisset quia multitudo maxima uenit ad eum dicit ad Philippum unde ememus panes ut manducent hii 6:6 hoc autem dicebat temptans eum 6:6 And this he said to prove him: for he himself knew

what he would do.

ipse enim sciebat quid esset facturus 6:7 respondit ei Phi- 6:7 Philip answered him, Two hundred pennyworth of

bread is not sufficient for them, that every one of them may take a little.

lippus ducentorum denariorum panes non sufficiunt eis ut

Biblia Sacra Vulgata

6:8 One of his disciples, Andrew, Simon Peter’s brother, unusquisque modicum quid accipiat 6:8 dicit ei unus ex di-

saith unto him,

6:9 There is a lad here, which hath five barley loaves, and scipulis eius Andreas frater Simonis Petri 6:9 est puer unus

two small fishes: but what are they among so many?

hic qui habet quinque panes hordiacios et duos pisces sed

6:10 And Jesus said, Make the men sit down. Now there haec quid sunt inter tantos 6:10 dixit ergo Iesus facite ho-

was much grass in the place. So the men sat down, in number about five thousand.

mines discumbere erat autem faenum multum in loco dis-

6:11 And Jesus took the loaves; and when he had given cubuerunt ergo uiri numero quasi quinque milia 6:11 acce-

thanks, he distributed to the disciples, and the disciples to them that were set down; and likewise of the fishes as

much as they would. pit ergo panes Iesus et cum gratias egisset distribuit discum-6:12 When they were filled, he said unto his disciples, bentibus similiter et ex piscibus quantum uolebant 6:12 ut

Gather up the fragments that remain, that nothing be lost.

autem impleti sunt dixit discipulis suis colligite quae supera-

6:13 Therefore they gathered them together, and filled uerunt fragmenta ne pereant 6:13 collegerunt ergo et imple-

twelve baskets with the fragments of the five barley loaves, which remained over and above unto them that had eaten.

uerunt duodecim cofinos fragmentorum ex quinque panibus

6:14 Then those men, when they had seen the miracle that hordiaciis quae superfuerunt his qui manducauerunt 6:14

Jesus did, said, This is of a truth that prophet that should come into the world.

illi ergo homines cum uidissent quod fecerat signum dicebant quia hic est uere propheta qui uenturus est in mundum

6:15 When Jesus therefore perceived that they would come 6:15 Iesus ergo cum cognouisset quia uenturi essent ut ra-

and take him by force, to make him a king, he departed again into a mountain himself alone.

perent eum et facerent eum regem fugit iterum in montem

6:16 And when even was now come, his disciples went ipse solus 6:16 ut autem sero factum est descenderunt dis-

down unto the sea, 6:17 And entered into a ship, and went over the sea toward cipuli eius ad mare 6:17 et cum ascendissent nauem uene-Capernaum. And it was now dark, and Jesus was not come to them.

runt trans mare in Capharnaum et tenebrae iam factae erant

6:18 And the sea arose by reason of a great wind that blew. et non uenerat ad eos Iesus 6:18 mare autem uento magno

6:19 So when they had rowed about five and twenty or flante exsurgebat 6:19 cum remigassent ergo quasi stadia

thirty furlongs, they see Jesus walking on the sea, and drawing nigh unto the ship: and they were afraid.

uiginti quinque aut triginta uident Iesum ambulantem super

6:20 But he saith unto them, It is I; be not afraid. mare et proximum naui fieri et timuerunt 6:20 ille autem 6:21 Then they willingly received him into the ship: and dicit eis ego sum nolite timere 6:21 uoluerunt ergo acci-

immediately the ship was at the land whither they went.

pere eum in naui et statim fuit nauis ad terram quam ibant

6:22 The day following, when the people which stood on 6:22 altera die turba quae stabat trans mare uidit quia na-

the other side of the sea saw that there was none other boat there, save that one whereinto his disciples were entered, and that Jesus went not with his disciples into the boat, but that his disciples were gone away alone;

uicula alia non erat ibi nisi una et quia non introisset cum discipulis suis Iesus in nauem sed soli discipuli eius abissent

6:23 (Howbeit there came other boats from Tiberias nigh 6:23 aliae uero superuenerunt naues a Tiberiade iuxta locum

unto the place where they did eat bread, after that the Lord had given thanks:)

ubi manducauerant panem gratias agente Domino 6:24 cum

6:24 When the people therefore saw that Jesus was not there, neither his disciples, they also took shipping, and came to Capernaum, seeking for Jesus.

ergo uidisset turba quia Iesus non esset ibi neque discipuli eius ascenderunt nauiculas et uenerunt Capharnaum quae-

6:25 And when they had found him on the other side of the rentes Iesum 6:25 et cum inuenissent eum trans mare dixe-

sea, they said unto him, Rabbi, when camest thou hither?

Euangelium Secundum Ioannem

runt ei rabbi quando huc uenisti 6:26 respondit eis Iesus et 6:26 Jesus answered them and said, Verily, verily, I say

unto you, Ye seek me, not because ye saw the miracles, but because ye did eat of the loaves, and were filled.

dixit amen amen dico uobis quaeritis me non quia uidistis signa sed quia manducastis ex panibus et saturati estis 6:27 6:27 Labour not for the meat which perisheth, but for that

meat which endureth unto everlasting life, which the Son of man shall give unto you: for him hath God the Father

operamini non cibum qui perit sed qui permanet in uitam sealed. aeternam quem Filius hominis uobis dabit hunc enim Pater signauit Deus 6:28 dixerunt ergo ad eum quid faciemus ut 6:28 Then said they unto him, What shall we do, that we

might work the works of God?

operemur opera Dei 6:29 respondit Iesus et dixit eis hoc est 6:29 Jesus answered and said unto them, This is the work

of God, that ye believe on him whom he hath sent.

opus Dei ut credatis in eum quem misit ille 6:30 dixerunt 6:30 They said therefore unto him, What sign shewest thou then, that we may see, and believe thee? what dost thou work?

ergo ei quod ergo tu facis signum ut uideamus et credamus tibi quid operaris 6:31 patres nostri manna manducauerunt 6:31 Our fathers did eat manna in the desert; as it is writ-

ten, He gave them bread from heaven to eat.

in deserto sicut scriptum est panem de caelo dedit eis man-

ducare 6:32 dixit ergo eis Iesus amen amen dico uobis non 6:32 Then Jesus said unto them, Verily, verily, I say unto you, Moses gave you not that bread from heaven; but my Father giveth you the true bread from heaven.

Moses dedit uobis panem de caelo sed Pater meus dat uobis panem de caelo uerum 6:33 panis enim Dei est qui descen- 6:33 For the bread of God is he which cometh down from

heaven, and giveth life unto the world.

dit de caelo et dat uitam mundo 6:34 dixerunt ergo ad eum 6:34 Then said they unto him, Lord, evermore give us this

bread.

Domine semper da nobis panem hunc 6:35 dixit autem eis 6:35 And Jesus said unto them, I am the bread of life: he

that cometh to me shall never hunger; and he that believeth on me shall never thirst.

Iesus ego sum panis uitae qui ueniet ad me non esuriet et qui credit in me non sitiet umquam 6:36 sed dixi uobis quia et 6:36 But I said unto you, That ye also have seen me, and

believe not.

uidistis me et non creditis 6:37 omne quod dat mihi Pater 6:37 All that the Father giveth me shall come to me; and

him that cometh to me I will in no wise cast out.

ad me ueniet et eum qui uenit ad me non eiciam foras 6:38 6:38 For I came down from heaven, not to do mine own

will, but the will of him that sent me.

quia descendi de caelo non ut faciam uoluntatem meam sed uoluntatem eius qui misit me 6:39 haec est autem uoluntas 6:39 And this is the Father’s will which hath sent me, that

of all which he hath given me I should lose nothing, but should raise it up again at the last day.

eius qui misit me Patris ut omne quod dedit mihi non perdam ex eo sed resuscitem illum nouissimo die 6:40 haec 6:40 And this is the will of him that sent me, that every

one which seeth the Son, and believeth on him, may have everlasting life: and I will raise him up at the last day.

est enim uoluntas Patris mei qui misit me ut omnis qui uidet Filium et credit in eum habeat uitam aeternam et resuscitabo ego eum in nouissimo die 6:41 murmurabant ergo Iudaei de 6:41 The Jews then murmured at him, because he said, I

am the bread which came down from heaven.

illo quia dixisset ego sum panis qui de caelo descendi 6:42 6:42 And they said, Is not this Jesus, the son of Joseph,

whose father and mother we know? how is it then that he saith, I came down from heaven?

et dicebant nonne hic est Iesus filius Ioseph cuius nos nouimus patrem et matrem quomodo ergo dicit hic quia de caelo descendi 6:43 respondit ergo Iesus et dixit eis nolite mur- 6:43 Jesus therefore answered and said unto them, Mur-

mur not among yourselves.

murare in inuicem 6:44 nemo potest uenire ad me nisi Pater 6:44 No man can come to me, except the Father which

hath sent me draw him: and I will raise him up at the last day.

qui misit me traxerit eum et ego resuscitabo eum nouissimo 6:45 It is written in the prophets, And they shall be all die 6:45 est scriptum in prophetis et erunt omnes docibiles

taught of God. Every man therefore that hath heard, and hath learned of the Father, cometh unto me.

Dei omnis qui audiuit a Patre et didicit uenit ad me 6:46

6:46 Not that any man hath seen the Father, save he which is of God, he hath seen the Father.

non quia Patrem uidit quisquam nisi is qui est a Deo hic ui-

6:47 Verily, verily, I say unto you, He that believeth on me dit Patrem 6:47 amen amen dico uobis qui credit in me hahath everlasting life.

6:48 I am that bread of life. bet uitam aeternam 6:48 ego sum panis uitae 6:49 patres

6:49 Your fathers did eat manna in the wilderness, and are dead.

uestri manducauerunt in deserto manna et mortui sunt 6:50

6:50 This is the bread which cometh down from heaven, that a man may eat thereof, and not die.

hic est panis de caelo descendens ut si quis ex ipso mandu-

6:51 I am the living bread which came down from heaven: cauerit non moriatur 6:51 ego sum panis uiuus qui de caelo

6:52 if any man eat of this bread, he shall live for ever: and descendi 6:52 si quis manducauerit ex hoc pane uiuet in the bread that I will give is my flesh, which I will give for the life of the world.

aeternum et panis quem ego dabo caro mea est pro mundi

6:53 The Jews therefore strove among themselves, saying, uita 6:53 litigabant ergo Iudaei ad inuicem dicentes quo-How can this man give us his flesh to eat?

modo potest hic nobis carnem suam dare ad manducandum

6:54 Then Jesus said unto them, Verily, verily, I say unto 6:54 dixit ergo eis Iesus amen amen dico uobis nisi manduyou, Except ye eat the flesh of the Son of man, and drink his blood, ye have no life in you.

caueritis carnem Filii hominis et biberitis eius sanguinem

6:55 Whoso eateth my flesh, and drinketh my blood, hath non habetis uitam in uobis 6:55 qui manducat meam careternal life; and I will raise him up at the last day.

nem et bibit meum sanguinem habet uitam aeternam et ego

6:56 For my flesh is meat indeed, and my blood is drink resuscitabo eum in nouissimo die 6:56 caro enim mea uere

indeed.

6:57 He that eateth my flesh, and drinketh my blood, est cibus et sanguis meus uere est potus 6:57 qui manducat

dwelleth in me, and I in him.

meam carnem et bibit meum sanguinem in me manet et ego

6:58 As the living Father hath sent me, and I live by the in illo 6:58 sicut misit me uiuens Pater et ego uiuo propter

Father: so he that eateth me, even he shall live by me.

6:59 This is that bread which came down from heaven: not Patrem et qui manducat me et ipse uiuet propter me 6:59 as your fathers did eat manna, and are dead: he that eateth of this bread shall live for ever.

hic est panis qui de caelo descendit non sicut manducauerunt patres uestri manna et mortui sunt qui manducat hunc pa-

6:60 These things said he in the synagogue, as he taught nem uiuet in aeternum 6:60 haec dixit in synagoga docens

in Capernaum.

6:61 Many therefore of his disciples, when they had heard in Capharnaum 6:61 multi ergo audientes ex discipulis eius

this, said, This is an hard saying; who can hear it?

6:62 When Jesus knew in himself that his disciples mur- dixerunt durus est hic sermo quis potest eum audire 6:62

mured at it, he said unto them, Doth this offend you?

sciens autem Iesus apud semet ipsum quia murmurarent de

6:63 What and if ye shall see the Son of man ascend up hoc discipuli eius dixit eis hoc uos scandalizat 6:63 si ergo

where he was before?

6:64 It is the spirit that quickeneth; the flesh profiteth noth- uideritis Filium hominis ascendentem ubi erat prius 6:64 ing: the words that I speak unto you, they are spirit, and they are life.

spiritus est qui uiuificat caro non prodeet quicquam uerba

6:65 But there are some of you that believe not. For Jesus quae ego locutus sum uobis spiritus et uita sunt 6:65 sed knew from the beginning who they were that believed not, and who should betray him.

sunt quidam ex uobis qui non credunt sciebat enim ab initio

6:66 And he said, Therefore said I unto you, that no man Iesus qui essent credentes et quis traditurus esset eum 6:66

can come unto me, except it were given unto him of my Father.

et dicebat propterea dixi uobis quia nemo potest uenire ad me nisi fuerit ei datum a Patre meo 6:67 ex hoc multi disci- 6:67 From that time many of his disciples went back, and

walked no more with him.

pulorum eius abierunt retro et iam non cum illo ambulabant 6:68 dixit ergo Iesus ad duodecim numquid et uos uultis abire 6:68 Then said Jesus unto the twelve, Will ye also go

away?

6:69 respondit ergo ei Simon Petrus Domine ad quem ibimus 6:69 Then Simon Peter answered him, Lord, to whom

shall we go? thou hast the words of eternal life.

uerba uitae aeternae habes 6:70 et nos credidimus et cog- 6:70 And we believe and are sure that thou art that Christ,

the Son of the living God.

nouimus quia tu es Christus Filius Dei 6:71 respondit eis 6:71 Jesus answered them, Have not I chosen you twelve,

and one of you is a devil?

Iesus nonne ego uos duodecim elegi et ex uobis unus diabolus est 6:72 dicebat autem Iudam Simonis Scariotis hic 6:72 He spake of Judas Iscariot the son of Simon: for he it

was that should betray him, being one of the twelve.

enim erat traditurus eum cum esset unus ex duodecim

7:1 post haec ambulabat Iesus in Galilaeam non enim uo- 7:1 After these things Jesus walked in Galilee: for he

would not walk in Jewry, because the Jews sought to kill him.

lebat in Iudaeam ambulare quia quaerebant eum Iudaei interficere 7:2 erat autem in proximo dies festus Iudaeorum sce- 7:2 Now the Jew’s feast of tabernacles was at hand. nopegia 7:3 dixerunt autem ad eum fratres eius transi hinc 7:3 His brethren therefore said unto him, Depart hence,

and go into Judaea, that thy disciples also may see the works that thou doest.

et uade in Iudaeam ut et discipuli tui uideant opera tua quae facis 7:4 nemo quippe in occulto quid facit et quaerit ipse 7:4 For there is no man that doeth any thing in secret, and

he himself seeketh to be known openly. If thou do these things, shew thyself to the world.

in palam esse si haec facis manifesta te ipsum mundo 7:5

7:5 For neither did his brethren believe in him.

neque enim fratres eius credebant in eum 7:6 dicit ergo eis 7:6 Then Jesus said unto them, My time is not yet come:

but your time is alway ready.

Iesus tempus meum nondum aduenit tempus autem uestrum

semper est paratum 7:7 non potest mundus odisse uos me 7:7 The world cannot hate you; but me it hateth, because I

testify of it, that the works thereof are evil.

autem odit quia ego testimonium perhibeo de illo quia opera

eius mala sunt 7:8 uos ascendite ad diem festum hunc ego 7:8 Go ye up unto this feast: I go not up yet unto this feast:

for my time is not yet full come.

non ascendo ad diem festum istum quia meum tempus nondum impletum est 7:9 haec cum dixisset ipse mansit in Ga- 7:9 When he had said these words unto them, he abode

still in Galilee.

lilaea 7:10 ut autem ascenderunt fratres eius tunc et ipse 7:10 But when his brethren were gone up, then went he

also up unto the feast, not openly, but as it were in secret.

ascendit ad diem festum non manifeste sed quasi in occulto 7:11 Iudaei ergo quaerebant eum in die festo et dicebant ubi 7:11 Then the Jews sought him at the feast, and said,

Where is he?

est ille 7:12 et murmur multus de eo erat in quidam enim di- 7:12 And there was much murmuring among the people

concerning him: for some said, He is a good man: others said, Nay; but he deceiveth the people.

cebant quia bonus est alii autem dicebant non sed seducit turbas 7:13 nemo tamen palam loquebatur de illo propter me- 7:13 Howbeit no man spake openly of him for fear of the

Jews.

tum Iudaeorum 7:14 iam autem die festo mediante ascen- 7:14 Now about the midst of the feast Jesus went up into

the temple, and taught.

dit Iesus in templum et docebat 7:15 et mirabantur Iudaei 7:15 And the Jews marvelled, saying, How knoweth this

man letters, having never learned?

dicentes quomodo hic litteras scit cum non didicerit 7:16 7:16 Jesus answered them, and said, My doctrine is not

mine, but his that sent me.

Biblia Sacra Vulgata

respondit eis Iesus et dixit mea doctrina non est mea sed

7:17 If any man will do his will, he shall know of the doc- eius qui misit me 7:17 si quis uoluerit uoluntatem eius fa-

trine, whether it be of God, or whether I speak of myself.

cere cognoscet de doctrina utrum ex Deo sit an ego a me

7:18 He that speaketh of himself seeketh his own glory: ipso loquar 7:18 qui a semet ipso loquitur gloriam prop-

but he that seeketh his glory that sent him, the same is true, and no unrighteousness is in him.

riam quaerit qui autem quaerit gloriam eius qui misit illum

7:19 Did not Moses give you the law, and yet none of you hic uerax est et iniustitia in illo non est 7:19 nonne Mo-

keepeth the law? Why go ye about to kill me?

7:20 The people answered and said, Thou hast a devil: ses dedit uobis legem et nemo ex uobis facit legem 7:20

who goeth about to kill thee?

quid me quaeritis interficere respondit turba et dixit daemo-

7:21 Jesus answered and said unto them, I have done one nium habes quis te quaerit interficere 7:21 respondit Iesus

work, and ye all marvel. 7:22 Moses therefore gave unto you circumcision; (not be- et dixit eis unum opus feci et omnes miramini 7:22 proptecause it is of Moses, but of the fathers;) and ye on the sabbath day circumcise a man.

rea Moses dedit uobis circumcisionem non quia ex Mose est

7:23 If a man on the sabbath day receive circumcision, sed ex patribus et in sabbato circumciditis hominem 7:23 si

that the law of Moses should not be broken; are ye angry at me, because I have made a man every whit whole on the

sabbath day? circumcisionem accipit homo in sabbato ut non soluatur lex

Mosi mihi indignamini quia totum hominem sanum feci in

7:24 Judge not according to the appearance, but judge sabbato 7:24 nolite iudicare secundum faciem sed iustum

righteous judgment.

7:25 Then said some of them of Jerusalem, Is not this he, iudicium iudicate 7:25 dicebant ergo quidam ex Hieroso-

whom they seek to kill? 7:26 But, lo, he speaketh boldly, and they say nothing unto lymis nonne hic est quem quaerunt interficere 7:26 et ecce him. Do the rulers know indeed that this is the very Christ?

palam loquitur et nihil ei dicunt numquid uere cognouerunt

7:27 Howbeit we know this man whence he is: but when principes quia hic est Christus 7:27 sed hunc scimus unde

Christ cometh, no man knoweth whence he is.

7:28 Then cried Jesus in the temple as he taught, saying, sit Christus autem cum uenerit nemo scit unde sit 7:28 cla-

Ye both know me, and ye know whence I am: and I am not come of myself, but he that sent me is true, whom ye

know not. mabat ergo docens in templo Iesus et dicens et me scitis et

unde sim scitis et a me ipso non ueni sed est uerus qui mi-

7:29 But I know him: for I am from him, and he hath sent sit me quem uos non scitis 7:29 ego scio eum quia ab ipso

me.

7:30 Then they sought to take him: but no man laid hands sum et ipse me misit 7:30 quaerebant ergo eum adprehen-

on him, because his hour was not yet come.

dere et nemo misit in illum manus quia nondum uenerat hora

7:31 And many of the people believed on him, and said, eius 7:31 de turba autem multi crediderunt in eum et dice-

When Christ cometh, will he do more miracles than these which this man hath done?

bant Christus cum uenerit numquid plura signa faciet quam

7:32 The Pharisees heard that the people murmured such quae hic facit 7:32 audierunt Pharisaei turbam murmuran-

things concerning him; and the Pharisees and the chief priests sent officers to take him.

tem de illo haec et miserunt principes et Pharisaei ministros

7:33 Then said Jesus unto them, Yet a little while am I ut adprehenderent eum 7:33 dixit ergo Iesus adhuc modi-

with you, and then I go unto him that sent me.

cum tempus uobiscum sum et uado ad eum qui misit me

7:34 Ye shall seek me, and shall not find me: and where I 7:34 quaeretis me et non inuenietis et ubi sum ego uos non

am, thither ye cannot come.

7:35 Then said the Jews among themselves, Whither will potestis uenire 7:35 dixerunt ergo Iudaei ad se ipsos quo

he go, that we shall not find him? will he go unto the dispersed among the Gentiles, and teach the Gentiles?

Euangelium Secundum Ioannem

hic iturus est quia non inueniemus eum numquid in dispersionem gentium iturus est et docturus gentes 7:36 quis est 7:36 What manner of saying is this that he said, Ye shall

seek me, and shall not find me: and where I am, thither ye cannot come?

hic sermo quem dixit quaeretis me et non inuenietis et ubi

sum ego non potestis uenire 7:37 in nouissimo autem die 7:37 In the last day, that great day of the feast, Jesus stood and cried, saying, If any man thirst, let him come unto me, and drink.

magno festiuitatis stabat Iesus et clamabat dicens si quis sitit ueniat ad me et bibat 7:38 qui credit in me sicut dixit sc- 7:38 He that believeth on me, as the scripture hath said,

out of his belly shall flow rivers of living water.

riptura flumina de uentre eius fluent aquae uiuae 7:39 hoc 7:39 (But this spake he of the Spirit, which they that be-

lieve on him should receive: for the Holy Ghost was not yet given; because that Jesus was not yet glorified.)

autem dixit de Spiritu quem accepturi erant credentes in eum non enim erat Spiritus quia Iesus nondum fuerat glorificatus 7:40 ex illa ergo turba cum audissent hos sermones eius dice- 7:40 Many of the people therefore, when they heard this

saying, said, Of a truth this is the Prophet.

bant hic est uere propheta 7:41 alii dicebant hic est Christus 7:41 Others said, This is the Christ. But some said, Shall

Christ come out of Galilee?

quidam autem dicebant numquid a Galilaea Christus uenit 7:42 nonne scriptura dicit quia ex semine Dauid et Bethleem 7:42 Hath not the scripture said, That Christ cometh of the

seed of David, and out of the town of Bethlehem, where David was?

castello ubi erat Dauid uenit Christus 7:43 dissensio itaque

7:43 So there was a division among the people because of him.

facta est in turba propter eum 7:44 quidam autem ex ip- 7:44 And some of them would have taken him; but no man

laid hands on him.

sis uolebant adprehendere eum sed nemo misit super illum manus 7:45 uenerunt ergo ministri ad pontifices et Phari- 7:45 Then came the officers to the chief priests and Phar-

isees; and they said unto them, Why have ye not brought him?

saeos et dixerunt eis illi quare non adduxistis eum 7:46 re-

7:46 The officers answered, Never man spake like this man.

sponderunt ministri numquam sic locutus est homo sicut hic homo 7:47 responderunt ergo eis Pharisaei numquid et uos 7:47 Then answered them the Pharisees, Are ye also de-

ceived?

seducti estis 7:48 numquid aliquis ex principibus credidit in 7:48 Have any of the rulers or of the Pharisees believed on

him?

eum aut ex Pharisaeis 7:49 sed turba haec quae non nouit 7:49 But this people who knoweth not the law are cursed. legem maledicti sunt 7:50 dicit Nicodemus ad eos ille qui 7:50 Nicodemus saith unto them, (he that came to Jesus

by night, being one of them,)

uenit ad eum nocte qui unus erat ex ipsis 7:51 numquid lex 7:51 Doth our law judge any man, before it hear him, and

know what he doeth?

nostra iudicat hominem nisi audierit ab ipso prius et cognouerit quid faciat 7:52 responderunt et dixerunt ei numquid 7:52 They answered and said unto him, Art thou also of

Galilee? Search, and look: for out of Galilee ariseth no prophet.

et tu Galilaeus es scrutare et uide quia propheta a Galilaea non surgit 7:53 et reuersi sunt unusquisque in domum suam 7:53 And every man went unto his own house.

8:1 Iesus autem perrexit in montem Oliueti 8:2 et dilu- 8:1 Jesus went unto the mount of Olives.

8:2 And early in the morning he came again into the temple, and all the people came unto him; and he sat down, and taught them.

culo iterum uenit in templum et omnis populus uenit ad eum et sedens docebat eos 8:3 adducunt autem scribae et Pha- 8:3 And the scribes and Pharisees brought unto him a

woman taken in adultery; and when they had set her in the midst,

risaei mulierem in adulterio deprehensam et statuerunt eam in medio 8:4 et dixerunt ei magister haec mulier modo de- 8:4 They say unto him, Master, this woman was taken in

adultery, in the very act.

Biblia Sacra Vulgata

8:5 Now Moses in the law commanded us, that such prehensa est in adulterio 8:5 in lege autem Moses mandauit

should be stoned: but what sayest thou?

8:6 This they said, tempting him, that they might have to nobis huiusmodi lapidare tu ergo quid dicis 8:6 haec autem

accuse him. But Jesus stooped down, and with his finger wrote on the ground, as though he heard them not.

dicebant temptantes eum ut possent accusare eum Iesus au-

8:7 So when they continued asking him, he lifted up him- tem inclinans se deorsum digito scribebat in terra 8:7 cum

self, and said unto them, He that is without sin among you, let him first cast a stone at her.

autem perseuerarent interrogantes eum erexit se et dixit eis qui sine peccato est uestrum primus in illam lapidem mittat

8:8 And again he stooped down, and wrote on the ground. 8:8 et iterum se inclinans scribebat in terra 8:9 audientes

8:9 And they which heard it, being convicted by their own conscience, went out one by one, beginning at the eldest, even unto the last: and Jesus was left alone, and the woman standing in the midst.

autem unus post unum exiebant incipientes a senioribus et

8:10 When Jesus had lifted up himself, and saw none but remansit solus et mulier in medio stans 8:10 erigens autem

the woman, he said unto her, Woman, where are those thine accusers? hath no man condemned thee?

se Iesus dixit ei mulier ubi sunt nemo te condemnauit 8:11

8:11 She said, No man, Lord. And Jesus said unto her, Neither do I condemn thee: go, and sin no more.

quae dixit nemo Domine dixit autem Iesus nec ego te con-

8:12 Then spake Jesus again unto them, saying, I am the demnabo uade et amplius iam noli peccare 8:12 iterum ergo

light of the world: he that followeth me shall not walk in darkness, but shall have the light of life.

locutus est eis Iesus dicens ego sum lux mundi qui sequitur

8:13 The Pharisees therefore said unto him, Thou bearest me non ambulabit in tenebris sed habebit lucem uitae 8:13 record of thyself; thy record is not true.

dixerunt ergo ei Pharisaei tu de te ipso testimonium perhi-

8:14 Jesus answered and said unto them, Though I bear bes testimonium tuum non est uerum 8:14 respondit Iesus

record of myself, yet my record is true: for I know whence I came, and whither I go; but ye cannot tell whence I come,

and whither I go. et dixit eis et si ego testimonium perhibeo de me ipso ue-

rum est testimonium meum quia scio unde ueni et quo uado

8:15 Ye judge after the flesh; I judge no man. uos autem nescitis unde uenio aut quo uado 8:15 uos se-

8:16 And yet if I judge, my judgment is true: for I am not cundum carnem iudicatis ego non iudico quemquam 8:16 alone, but I and the Father that sent me.

et si iudico ego iudicium meum uerum est quia solus non

8:17 It is also written in your law, that the testimony of sum sed ego et qui me misit Pater 8:17 et in lege uestra two men is true.

scriptum est quia duorum hominum testimonium uerum est

8:18 I am one that bear witness of myself, and the Father 8:18 ego sum qui testimonium perhibeo de me ipso et testi-

that sent me beareth witness of me. 8:19 Then said they unto him, Where is thy Father? Jesus monium perhibet de me qui misit me Pater 8:19 dicebant answered, Ye neither know me, nor my Father: if ye had known me, ye should have known my Father also.

ergo ei ubi est Pater tuus respondit Iesus neque me scitis neque Patrem meum si me sciretis forsitan et Patrem meum

8:20 These words spake Jesus in the treasury, as he taught sciretis 8:20 haec uerba locutus est in gazofilacio docens in

in the temple: and no man laid hands on him; for his hour was not yet come.

templo et nemo adprehendit eum quia necdum uenerat hora

8:21 Then said Jesus again unto them, I go my way, and eius 8:21 dixit ergo iterum eis Iesus ego uado et quaeretis

ye shall seek me, and shall die in your sins: whither I go, ye cannot come.

me et in peccato uestro moriemini quo ego uado uos non po-

8:22 Then said the Jews, Will he kill himself? because he testis uenire 8:22 dicebant ergo Iudaei numquid interficiet

saith, Whither I go, ye cannot come.

semet ipsum quia dicit quo ego uado uos non potestis uenire

Euangelium Secundum Ioannem

8:23 et dicebat eis uos de deorsum estis ego de supernis sum 8:23 And he said unto them, Ye are from beneath; I am

from above: ye are of this world; I am not of this world.

uos de mundo hoc estis ego non sum de hoc mundo 8:24 8:24 I said therefore unto you, that ye shall die in your sins: for if ye believe not that I am he, ye shall die in your sins.

dixi ergo uobis quia moriemini in peccatis uestris si enim non credideritis quia ego sum moriemini in peccato uestro 8:25 dicebant ergo ei tu quis es dixit eis Iesus principium quia 8:25 Then said they unto him, Who art thou? And Jesus

saith unto them, Even the same that I said unto you from the beginning.

et loquor uobis 8:26 multa habeo de uobis loqui et iudicare

8:26 I have many things to say and to judge of you: but he that sent me is true; and I speak to the world those things which I have heard of him.

sed qui misit me uerax est et ego quae audiui ab eo haec loquor in mundo 8:27 et non cognouerunt quia Patrem eis 8:27 They understood not that he spake to them of the Fa-

ther.

dicebat 8:28 dixit ergo eis Iesus cum exaltaueritis Filium 8:28 Then said Jesus unto them, When ye have lifted up

the Son of man, then shall ye know that I am he, and that I do nothing of myself; but as my Father hath taught me, I

hominis tunc cognoscetis quia ego sum et a me ipso facio speak these things. nihil sed sicut docuit me Pater haec loquor 8:29 et qui me 8:29 And he that sent me is with me: the Father hath not

left me alone; for I do always those things that please him.

misit mecum est non reliquit me solum quia ego quae placita sunt ei facio semper 8:30 haec illo loquente multi credide- 8:30 As he spake these words, many believed on him. runt in eum 8:31 dicebat ergo Iesus ad eos qui crediderunt 8:31 Then said Jesus to those Jews which believed on him,

If ye continue in my word, then are ye my disciples indeed;

ei Iudaeos si uos manseritis in sermone meo uere discipuli mei eritis 8:32 et cognoscetis ueritatem et ueritas liberabit 8:32 And ye shall know the truth, and the truth shall make

you free.

uos 8:33 responderunt ei semen Abrahae sumus et nemini 8:33 They answered him, We be Abraham’s seed, and

were never in bondage to any man: how sayest thou, Ye shall be made free?

seruiuimus umquam quomodo tu dicis liberi eritis 8:34 re-

8:34 Jesus answered them, Verily, verily, I say unto you, Whosoever committeth sin is the servant of sin.

spondit eis Iesus amen amen dico uobis quia omnis qui facit

peccatum seruus est peccati 8:35 seruus autem non manet 8:35 And the servant abideth not in the house for ever: but

the Son abideth ever.

in domo in aeternum filius manet in aeternum 8:36 si ergo 8:36 If the Son therefore shall make you free, ye shall be

free indeed.

Filius uos liberauerit uere liberi eritis 8:37 scio quia filii 8:37 I know that ye are Abraham’s seed; but ye seek to kill

me, because my word hath no place in you.

Abrahae estis sed quaeritis me interficere quia sermo meus non capit in uobis 8:38 ego quod uidi apud Patrem loquor 8:38 I speak that which I have seen with my Father: and

ye do that which ye have seen with your father.

et uos quae uidistis apud patrem uestrum facitis 8:39 res- 8:39 They answered and said unto him, Abraham is our

father. Jesus saith unto them, If ye were Abraham’s children, ye would do the works of Abraham.

ponderunt et dixerunt ei pater noster Abraham est dicit eis Iesus si filii Abrahae estis opera Abrahae facite 8:40 nunc 8:40 But now ye seek to kill me, a man that hath told you

the truth, which I have heard of God: this did not Abraham.

autem quaeritis me interficere hominem qui ueritatem uobis locutus sum quam audiui a Deo hoc Abraham non fecit 8:41 uos facitis opera patris uestri dixerunt itaque ei nos ex 8:41 Ye do the deeds of your father. Then said they to him,

We be not born of fornication; we have one Father, even God.

fornicatione non sumus nati unum patrem habemus Deum 8:42 dixit ergo eis Iesus si Deus pater uester esset diligeretis 8:42 Jesus said unto them, If God were your Father, ye

would love me: for I proceeded forth and came from God; neither came I of myself, but he sent me.

utique me ego enim ex Deo processi et ueni neque enim a

Biblia Sacra Vulgata

8:43 Why do ye not understand my speech? even because me ipso ueni sed ille me misit 8:43 quare loquellam meam

ye cannot hear my word.

non cognoscitis quia non potestis audire sermonem meum

8:44 Ye are of your father the devil, and the lusts of your 8:44 uos ex patre diabolo estis et desideria patris uestri uul-

father ye will do. He was a murderer from the beginning, and abode not in the truth, because there is no truth in him. When he speaketh a lie, he speaketh of his own: for he is a liar, and the father of it.

tis facere ille homicida erat ab initio et in ueritate non stetit quia non est ueritas in eo cum loquitur mendacium ex pro-

8:45 And because I tell you the truth, ye believe me not. priis loquitur quia mendax est et pater eius 8:45 ego autem 8:46 Which of you convinceth me of sin? And if I say the quia ueritatem dico non creditis mihi 8:46 quis ex uobis ar-

truth, why do ye not believe me?

guit me de peccato si ueritatem dico quare uos non credi-

8:47 He that is of God heareth God’s words: ye therefore tis mihi 8:47 qui est ex Deo uerba Dei audit propterea uos

hear them not, because ye are not of God.

8:48 Then answered the Jews, and said unto him, Say we non auditis quia ex Deo non estis 8:48 responderunt igitur

not well that thou art a Samaritan, and hast a devil?

Iudaei et dixerunt ei nonne bene dicimus nos quia Samari-

8:49 Jesus answered, I have not a devil; but I honour my tanus es tu et daemonium habes 8:49 respondit Iesus ego

Father, and ye do dishonour me.

daemonium non habeo sed honorifico Patrem meum et uos

8:50 And I seek not mine own glory: there is one that inhonoratis me 8:50 ego autem non quaero gloriam meam

seeketh and judgeth.

8:51 Verily, verily, I say unto you, If a man keep my say- est qui quaerit et iudicat 8:51 amen amen dico uobis si quis

ing, he shall never see death.

sermonem meum seruauerit mortem non uidebit in aeternum

8:52 Then said the Jews unto him, Now we know that thou 8:52 dixerunt ergo Iudaei nunc cognouimus quia daemonium

hast a devil. Abraham is dead, and the prophets; and thou sayest, If a man keep my saying, he shall never taste of

death. habes Abraham mortuus est et prophetae et tu dicis si quis

sermonem meum seruauerit non gustabit mortem in aeter-

8:53 Art thou greater than our father Abraham, which is num 8:53 numquid tu maior es patre nostro Abraham qui dead? and the prophets are dead: whom makest thou thyself?

mortuus est et prophetae mortui sunt quem te ipsum facis

8:54 Jesus answered, If I honour myself, my honour is 8:54 respondit Iesus si ego glorifico me ipsum gloria mea ni-

nothing: it is my Father that honoureth me; of whom ye say, that he is your God:

hil est est Pater meus qui glorificat me quem uos dicitis quia

8:55 Yet ye have not known him; but I know him: and if I Deus noster est 8:55 et non cognouistis eum ego autem noui

should say, I know him not, I shall be a liar like unto you: but I know him, and keep his saying.

eum et si dixero quia non scio eum ero similis uobis mendax

8:56 Your father Abraham rejoiced to see my day: and he sed scio eum et sermonem eius seruo 8:56 Abraham pasaw it, and was glad.

ter uester exultauit ut uideret diem meum et uidit et gauisus

8:57 Then said the Jews unto him, Thou art not yet fifty est 8:57 dixerunt ergo Iudaei ad eum quinquaginta annos

years old, and hast thou seen Abraham? 8:58 Jesus said unto them, Verily, verily, I say unto you, nondum habes et Abraham uidisti 8:58 dixit eis Iesus amen Before Abraham was, I am.

8:59 Then took they up stones to cast at him: but Jesus amen dico uobis antequam Abraham fieret ego sum 8:59 tu-

hid himself, and went out of the temple, going through the midst of them, and so passed by.

lerunt ergo lapides ut iacerent in eum Iesus autem abscondit se et exiuit de templo

9:1 And as Jesus passed by, he saw a man which was blind 9:1 et praeteriens uidit hominem caecum a natiuitate 9:2

from his birth. 9:2 And his disciples asked him, saying, Master, who did sin, this man, or his parents, that he was born blind?

Euangelium Secundum Ioannem

et interrogauerunt eum discipuli sui rabbi quis peccauit hic aut parentes eius ut caecus nasceretur 9:3 respondit Iesus 9:3 Jesus answered, Neither hath this man sinned, nor his

parents: but that the works of God should be made manifest in him.

neque hic peccauit neque parentes eius sed ut manifestetur opera Dei in illo 9:4 me oportet operari opera eius qui mi- 9:4 I must work the works of him that sent me, while it is

day: the night cometh, when no man can work.

sit me donec dies est uenit nox quando nemo potest operari 9:5 quamdiu in mundo sum lux sum mundi 9:6 haec cum 9:5 As long as I am in the world, I am the light of the

world.

9:6 When he had thus spoken, he spat on the ground, and made clay of the spittle, and he anointed the eyes of the blind man with the clay,

dixisset expuit in terram et fecit lutum ex sputo et linuit lutum super oculos eius 9:7 et dixit ei uade laua in natato- 9:7 And said unto him, Go, wash in the pool of Siloam,

(which is by interpretation, Sent.) He went his way therefore, and washed, and came seeing.

ria Siloae quod interpretatur Missus abiit ergo et lauit et uenit uidens 9:8 itaque uicini et qui uidebant eum prius quia 9:8 The neighbours therefore, and they which before had

seen him that he was blind, said, Is not this he that sat and begged?

mendicus erat dicebant nonne hic est qui sedebat et mendicabat alii dicebant quia hic est 9:9 alii autem nequaquam 9:9 Some said, This is he: others said, He is like him: but

he said, I am he.

sed similis est eius ille dicebat quia ego sum 9:10 dicebant 9:10 Therefore said they unto him, How were thine eyes

opened?

ergo ei quomodo aperti sunt oculi tibi 9:11 respondit ille 9:11 He answered and said, A man that is called Jesus

made clay, and anointed mine eyes, and said unto me, Go to the pool of Siloam, and wash: and I went and washed,

homo qui dicitur Iesus lutum fecit et unxit oculos meos et and I received sight. dixit mihi uade ad natatoriam Siloae et laua et abii et laui et uidi 9:12 dixerunt ei ubi est ille ait nescio 9:13 adducunt 9:12 Then said they unto him, Where is he? He said, I

know not.

9:13 They brought to the Pharisees him that aforetime was blind.

eum ad Pharisaeos qui caecus fuerat 9:14 erat autem sab-

9:14 And it was the sabbath day when Jesus made the clay,

batum quando lutum fecit Iesus et aperuit oculos eius 9:15 ^{and opened his eyes.}

9:15 Then again the Pharisees also asked him how he had received his sight. He said unto them, He put clay upon mine eyes, and I washed, and do see.

iterum ergo interrogabant eum Pharisaei quomodo uidisset ille autem dixit eis lutum posuit mihi super oculos et laui et uideo 9:16 dicebant ergo ex Pharisaeis quidam non est hic 9:16 Therefore said some of the Pharisees, This man is not

of God, because he keepeth not the sabbath day. Others said, How can a man that is a sinner do such miracles?

homo a Deo quia sabbatum non custodit alii dicebant quo- And there was a division among them. modo potest homo peccator haec signa facere et scisma erat in eis 9:17 dicunt ergo caeco iterum tu quid dicis de eo qui 9:17 They say unto the blind man again, What sayest thou

of him, that he hath opened thine eyes? He said, He is a prophet.

aperuit oculos tuos ille autem dixit quia propheta est 9:18

9:18 But the Jews did not believe concerning him, that he had been blind, and received his sight, until they called the parents of him that had received his sight.

non crediderunt ergo Iudaei de illo quia caecus fuisset et uidisset donec uocauerunt parentes eius qui uiderat 9:19 et 9:19 And they asked them, saying, Is this your son, who

ye say was born blind? how then doth he now see?

interrogauerunt eos dicentes hic est filius uester quem uos dicitis quia caecus natus est quomodo ergo nunc uidet 9:20 9:20 His parents answered them and said, We know that

this is our son, and that he was born blind:

responderunt eis parentes eius et dixerunt scimus quia hic est filius noster et quia caecus natus est 9:21 quomodo autem 9:21 But by what means he now seeth, we know not; or

who hath opened his eyes, we know not: he is of age; ask him: he shall speak for himself.

nunc uideat nescimus aut quis eius aperuit oculos nos nes-

Biblia Sacra Vulgata

cimus ipsum interrogate aetatem habet ipse de se loquatur

9:22 These words spake his parents, because they feared 9:22 haec dixerunt parentes eius quia timebant Iudaeos iam

the Jews: for the Jews had agreed already, that if any man did confess that he was Christ, he should be put out of the

synagogue. enim conspirauerant Iudaei ut si quis eum confiteretur Ch-9:23 Therefore said his parents, He is of age; ask him. ristum extra synagogam fieret 9:23 propterea parentes eius 9:24 Then again called they the man that was blind, and dixerunt quia aetatem habet ipsum interrogate 9:24 uoca-

said unto him, Give God the praise: we know that this man is a sinner.

uerunt ergo rursum hominem qui fuerat caecus et dixerunt ei

9:25 He answered and said, Whether he be a sinner or no, da gloriam Deo nos scimus quia hic homo peccator est 9:25

I know not: one thing I know, that, whereas I was blind, now I see.

dixit ergo ille si peccator est nescio unum scio quia caecus

9:26 Then said they to him again, What did he to thee? cum essem modo uideo 9:26 dixerunt ergo illi quid fecit

how opened he thine eyes?

9:27 He answered them, I have told you already, and ye tibi quomodo aperuit tibi oculos 9:27 respondit eis dixi uo-

did not hear: wherefore would ye hear it again? will ye also be his disciples?

bis iam et audistis quid iterum uultis audire numquid et uos

9:28 Then they reviled him, and said, Thou art his disciple; uultis discipuli eius fieri 9:28 maledixerunt ei et dixerunt

but we are Moses’ disciples.

9:29 We know that God spake unto Moses: as for this fel- tu discipulus illius es nos autem Mosi discipuli sumus 9:29

low, we know not from whence he is.

nos scimus quia Mosi locutus est Deus hunc autem nesci-

9:30 The man answered and said unto them, Why herein mus unde sit 9:30 respondit ille homo et dixit eis in hoc

is a marvellous thing, that ye know not from whence he is, and yet he hath opened mine eyes.

enim mirabile est quia uos nescitis unde sit et aperuit meos

9:31 Now we know that God heareth not sinners: but if oculos 9:31 scimus autem quia peccatores Deus non audit

any man be a worshipper of God, and doeth his will, him he heareth.

sed si quis Dei cultor est et uoluntatem eius facit hunc exau-

9:32 Since the world began was it not heard that any man dit 9:32 a saeculo non est auditum quia aperuit quis oculos

opened the eyes of one that was born blind.

9:33 If this man were not of God, he could do nothing. caeci nati 9:33 nisi esset hic a Deo non poterat facere quic-9:34 They answered and said unto him, Thou wast alto- quam 9:34 responderunt et dixerunt ei in peccatis natus es

gether born in sins, and dost thou teach us? And they cast him out.

totus et tu doces nos et eiecerunt eum foras 9:35 audiuit

9:35 Jesus heard that they had cast him out; and when he had found him, he said unto him, Dost thou believe on the Son of God?

Iesus quia eiecerunt eum foras et cum inuenisset eum dixit

9:36 He answered and said, Who is he, Lord, that I might ei tu credis in Filium Dei 9:36 respondit ille et dixit quis

believe on him?

9:37 And Jesus said unto him, Thou hast both seen him, est Domine ut credam in eum 9:37 et dixit ei Iesus et ui-

and it is he that talketh with thee.

9:38 And he said, Lord, I believe. And he worshipped disti eum et qui loquitur tecum ipse est 9:38 at ille ait credo

him.

9:39 And Jesus said, For judgment I am come into this Domine et procidens adorauit eum 9:39 dixit ei Iesus in iu-

world, that they which see not might see; and that they which see might be made blind.

dicium ego in hunc mundum ueni ut qui non uident uideant

9:40 And some of the Pharisees which were with him et qui uident caeci fiant 9:40 et audierunt ex Pharisaeis qui

heard these words, and said unto him, Are we blind also?

cum ipso erant et dixerunt ei numquid et nos caeci sumus

9:41 Jesus said unto them, If ye were blind, ye should 9:41 dixit eis Iesus si caeci essetis non haberetis peccatum

have no sin: but now ye say, We see; therefore your sin remaineth.

nunc uero dicitis quia uidemus peccatum uestrum manet

10:1 Verily, verily, I say unto you, He that entereth not by 10:1 amen amen dico uobis qui non intrat per ostium in the door into the sheepfold, but climbeth up some other way, the same is a thief and a robber.

ouile ouium sed ascendit aliunde ille fur est et latro 10:2 10:2 But he that entereth in by the door is the shepherd of

the sheep.

qui autem intrat per ostium pastor est ouium 10:3 huic os- 10:3 To him the porter openeth; and the sheep hear his

voice: and he calleth his own sheep by name, and leadeth them out.

tiarius aperit et oues uocem eius audiunt et proprias oues uocat nominatim et educit eas 10:4 et cum proprias oues 10:4 And when he putteth forth his own sheep, he goeth

before them, and the sheep follow him: for they know his voice.

emiserit ante eas uadit et oues illum sequuntur quia sciunt

uocem eius 10:5 alienum autem non sequuntur sed fugient 10:5 And a stranger will they not follow, but will flee from

him: for they know not the voice of strangers.

ab eo quia non nouerunt uocem alienorum 10:6 hoc prouer- 10:6 This parable spake Jesus unto them: but they under-

stood not what things they were which he spake unto them.

bium dixit eis Iesus illi autem non cognouerunt quid loquere-

tur eis 10:7 dixit ergo eis iterum Iesus amen amen dico uo- 10:7 Then said Jesus unto them again, Verily, verily, I say

unto you, I am the door of the sheep.

bis quia ego sum ostium ouium 10:8 omnes quotquot uene- 10:8 All that ever came before me are thieves and robbers:

but the sheep did not hear them.

runt fures sunt et latrones sed non audierunt eos oues 10:9 10:9 I am the door: by me if any man enter in, he shall be

saved, and shall go in and out, and find pasture.

ego sum ostium per me si quis introierit saluabitur et ingredietur et egredietur et pascua inueniet 10:10 fur non uenit 10:10 The thief cometh not, but for to steal, and to kill,

and to destroy: I am come that they might have life, and that they might have it more abundantly.

nisi ut furetur et mactet et perdat ego ueni ut uitam habeant et abundantius habeant 10:11 ego sum pastor bonus bonus 10:11 I am the good shepherd: the good shepherd giveth

his life for the sheep.

pastor animam suam dat pro ouibus 10:12 mercennarius et 10:12 But he that is an hireling, and not the shepherd,

whose own the sheep are not, seeth the wolf coming, and leaveth the sheep, and fleeth: and the wolf catcheth them,

qui non est pastor cuius non sunt oues propriae uidet lupum and scattereth the sheep. uenientem et dimittit oues et fugit et lupus rapit et dispergit oues 10:13 mercennarius autem fugit quia mercennarius 10:13 The hireling fleeth, because he is an hireling, and

careth not for the sheep.

est et non pertinet ad eum de ouibus 10:14 ego sum pastor 10:14 I am the good shepherd, and know my sheep, and

am known of mine.

bonus et cognosco meas et cognoscunt me meae 10:15 sic- 10:15 As the Father knoweth me, even so know I the Fa-

ther: and I lay down my life for the sheep.

ut nouit me Pater et ego agnosco Patrem et animam meam

pono pro ouibus 10:16 et alias oues habeo quae non sunt 10:16 And other sheep I have, which are not of this fold: them also I must bring, and they shall hear my voice; and there shall be one fold, and one shepherd.

ex hoc ouili et illas oportet me adducere et uocem meam audient et fiet unum ouile unus pastor 10:17 propterea me 10:17 Therefore doth my Father love me, because I lay

down my life, that I might take it again.

Pater diligit quia ego pono animam meam ut iterum sumam

eam 10:18 nemo tollit eam a me sed ego pono eam a me 10:18 No man taketh it from me, but I lay it down of myself. I have power to lay it down, and I have power to take it again. This commandment have I received of my Father.

ipso potestatem habeo ponendi eam et potestatem habeo iterum sumendi eam hoc mandatum accepi a Patre meo 10:19 10:19 There was a division therefore again among the

Jews for these sayings.

dissensio iterum facta est inter Iudaeos propter sermones hos 10:20 dicebant autem multi ex ipsis daemonium habet et insa- 10:20 And many of them said, He hath a devil, and is mad;

why hear ye him?

nit quid eum auditis 10:21 alii dicebant haec uerba non sunt 10:21 Others said, These are not the words of him that hath

a devil. Can a devil open the eyes of the blind?

daemonium habentis numquid daemonium potest caecorum

Biblia Sacra Vulgata

10:22 And it was at Jerusalem the feast of the dedication, oculos aperire 10:22 facta sunt autem encenia in Hieroso-

and it was winter.

10:23 And Jesus walked in the temple in Solomon’s porch. lymis et hiemps erat 10:23 et ambulabat Iesus in templo in 10:24 Then came the Jews round about him, and said unto porticu Salomonis 10:24 circumdederunt ergo eum Iudaei

him, How long dost thou make us to doubt? If thou be the Christ, tell us plainly.

et dicebant ei quousque animam nostram tollis si tu es Chri-

10:25 Jesus answered them, I told you, and ye believed stus dic nobis palam 10:25 respondit eis Iesus loquor uobis

not: the works that I do in my Father’s name, they bear witness of me.

et non creditis opera quae ego facio in nomine Patris mei

10:26 But ye believe not, because ye are not of my sheep, haec testimonium perhibent de me 10:26 sed uos non creas I said unto you.

10:27 My sheep hear my voice, and I know them, and they ditis quia non estis ex ouibus meis 10:27 oues meae uocem

follow me: 10:28 And I give unto them eternal life; and they shall meam audiunt et ego cognosco eas et sequuntur me 10:28 never perish, neither shall any man pluck them out of my hand.

et ego uitam aeternam do eis et non peribunt in aeternum

10:29 My Father, which gave them me, is greater than all; et non rapiet eas quisquam de manu mea 10:29 Pater meus and no man is able to pluck them out of my Father’s hand.

quod dedit mihi maius omnibus est et nemo potest rapere de

10:30 I and my Father are one. manu Patris mei 10:30 ego et Pater unum sumus 10:31 su-

10:31 Then the Jews took up stones again to stone him.

10:32 Jesus answered them, Many good works have I stulerunt lapides Iudaei ut lapidarent eum 10:32 respondit

shewed you from my Father; for which of those works do ye stone me?

eis Iesus multa opera bona ostendi uobis ex Patre meo pro-

10:33 The Jews answered him, saying, For a good work pter quod eorum opus me lapidatis 10:33 responderunt ei

we stone thee not; but for blasphemy; and because that thou, being a man, makest thyself God.

Iudaei de bono opere non lapidamus te sed de blasphemia et

10:34 Jesus answered them, Is it not written in your law, I quia tu homo cum sis facis te ipsum Deum 10:34 respondit said, Ye are gods?

eis Iesus nonne scriptum est in lege uestra quia ego dixi dii

10:35 If he called them gods, unto whom the word of God estis 10:35 si illos dixit deos ad quos sermo Dei factus est et

came, and the scripture cannot be broken;

10:36 Say ye of him, whom the Father hath sanctified, and non potest solui scriptura 10:36 quem Pater sanctificauit et

sent into the world, Thou blasphemest; because I said, I am the Son of God?

misit in mundum uos dicitis quia blasphemas quia dixi Filius

10:37 If I do not the works of my Father, believe me not. Dei sum 10:37 si non facio opera Patris mei nolite credere 10:38 But if I do, though ye believe not me, believe the mihi 10:38 si autem facio et si mihi non uultis credere ope-

works: that ye may know, and believe, that the Father is in me, and I in him.

ribus credite ut cognoscatis et credatis quia in me est Pater

10:39 Therefore they sought again to take him: but he es- et ego in Patre 10:39 quaerebant ergo eum prendere et excaped out of their hand,

10:40 And went away again beyond Jordan into the place iuit de manibus eorum 10:40 et abiit iterum trans Iordanen

where John at first baptized; and there he abode.

in eum locum ubi erat Iohannes baptizans primum et mansit

10:41 And many resorted unto him, and said, John did no illic 10:41 et multi uenerunt ad eum et dicebant quia Iohan-

miracle: but all things that John spake of this man were true.

nes quidem signum fecit nullum 10:42 omnia autem quae-

10:42 And many believed on him there.

cumque dixit Iohannes de hoc uera erant et multi crediderunt in eum

11:1 Now a certain man was sick, named Lazarus, of 11:1 erat autem quidam languens Lazarus a Bethania de

Bethany, the town of Mary and her sister Martha.

castello Mariae et Marthae sororis eius 11:2 Maria autem 11:2 (It was that Mary which anointed the Lord with oint-

ment, and wiped his feet with her hair, whose brother Lazarus was sick.)

erat quae unxit Dominum unguento et extersit pedes eius capillis suis cuius frater Lazarus infirmabatur 11:3 miserunt 11:3 Therefore his sisters sent unto him, saying, Lord, be-

hold, he whom thou lovest is sick.

ergo sorores ad eum dicentes Domine ecce quem amas infirmatur 11:4 audiens autem Iesus dixit eis infirmitas haec 11:4 When Jesus heard that, he said, This sickness is not

unto death, but for the glory of God, that the Son of God might be glorified thereby.

non est ad mortem sed pro gloria Dei ut glorificetur Filius

Dei per eam 11:5 diligebat autem Iesus Martham et soro- 11:5 Now Jesus loved Martha, and her sister, and Lazarus.

rem eius Mariam et Lazarum 11:6 ut ergo audiuit quia in- 11:6 When he had heard therefore that he was sick, he

abode two days still in the same place where he was.

firmabatur tunc quidem mansit in eodem loco duobus diebus 11:7 deinde post haec dicit discipulis suis eamus in Iu- 11:7 Then after that saith he to his disciples, Let us go into

Judaea again.

daeam iterum 11:8 dicunt ei discipuli rabbi nunc quaere- 11:8 His disciples say unto him, Master, the Jews of late

sought to stone thee; and goest thou thither again?

bant te Iudaei lapidare et iterum uadis illuc 11:9 respondit 11:9 Jesus answered, Are there not twelve hours in the

day? If any man walk in the day, he stumbleth not, because he seeth the light of this world.

Iesus nonne duodecim horae sunt diei si quis ambulauerit in die non offendit quia lucem huius mundi uidet 11:10 si au- 11:10 But if a man walk in the night, he stumbleth, be-

cause there is no light in him.

tem ambulauerit nocte offendit quia lux non est in eo 11:11 11:11 These things said he: and after that he saith unto

them, Our friend Lazarus sleepeth; but I go, that I may awake him out of sleep.

haec ait et post hoc dicit eis Lazarus amicus noster dormit

sed uado ut a somno exsuscitem eum 11:12 dixerunt ergo 11:12 Then said his disciples, Lord, if he sleep, he shall

do well.

discipuli eius Domine si dormit saluus erit 11:13 dixerat au- 11:13 Howbeit Jesus spake of his death: but they thought

that he had spoken of taking of rest in sleep.

tem Iesus de morte eius illi autem putauerunt quia de dormitione somni diceret 11:14 tunc ergo dixit eis Iesus manifeste 11:14 Then said Jesus unto them plainly, Lazarus is dead. Lazarus mortuus est 11:15 et gaudeo propter uos ut creda- 11:15 And I am glad for your sakes that I was not there, to

the intent ye may believe; nevertheless let us go unto him.

tis quoniam non eram ibi sed eamus ad eum 11:16 dixit 11:16 Then said Thomas, which is called Didymus, unto his fellowdisciples, Let us also go, that we may die with him.

ergo Thomas qui dicitur Didymus ad condiscipulos eamus et nos ut moriamur cum eo 11:17 uenit itaque Iesus et inuenit 11:17 Then when Jesus came, he found that he had lain in

the grave four days already.

eum quattuor dies iam in monumento habentem 11:18 erat 11:18 Now Bethany was nigh unto Jerusalem, about fif-

teen furlongs off:

autem Bethania iuxta Hierosolyma quasi stadiis quindecim 11:19 multi autem ex Iudaeis uenerant ad Martham et Mariam 11:19 And many of the Jews came to Martha and Mary, to

comfort them concerning their brother.

ut consolarentur eas de fratre suo 11:20 Martha ergo ut au- 11:20 Then Martha, as soon as she heard that Jesus was

coming, went and met him: but Mary sat still in the house.

diuit quia Iesus uenit occurrit illi Maria autem domi sedebat 11:21 dixit ergo Martha ad Iesum Domine si fuisses hic frater 11:21 Then said Martha unto Jesus, Lord, if thou hadst

been here, my brother had not died.

meus non fuisset mortuus 11:22 sed et nunc scio quia quae- 11:22 But I know, that even now, whatsoever thou wilt ask

of God, God will give it thee.

cumque poposceris a Deo dabit tibi Deus 11:23 dicit illi Ie- 11:23 Jesus saith unto her, Thy brother shall rise again. sus resurget frater tuus 11:24 dicit ei Martha scio quia re- 11:24 Martha saith unto him, I know that he shall rise

again in the resurrection at the last day.

Biblia Sacra Vulgata

11:25 Jesus said unto her, I am the resurrection, and the surget in resurrectione in nouissima die 11:25 dixit ei Iesus

life: he that believeth in me, though he were dead, yet shall he live:

ego sum resurrectio et uita qui credit in me et si mortuus fue-

11:26 And whosoever liveth and believeth in me shall rit uiuet 11:26 et omnis qui uiuit et credit in me non morie-

never die. Believest thou this?

11:27 She saith unto him, Yea, Lord: I believe that thou tur in aeternum credis hoc 11:27 ait illi utique Domine ego

art the Christ, the Son of God, which should come into the world.

credidi quia tu es Christus Filius Dei qui in mundum uenisti

11:28 And when she had so said, she went her way, and 11:28 et cum haec dixisset abiit et uocauit Mariam sororem

called Mary her sister secretly, saying, The Master is come, and calleth for thee.

suam silentio dicens magister adest et uocat te 11:29 illa

11:29 As soon as she heard that, she arose quickly, and came unto him.

11:30 Now Jesus was not yet come into the town, but was ut audiuit surgit cito et uenit ad eum 11:30 nondum enim

in that place where Martha met him.

uenerat Iesus in castellum sed erat adhuc in illo loco ubi oc-

11:31 The Jews then which were with her in the house, currerat ei Martha 11:31 Iudaei igitur qui erant cum ea in

and comforted her, when they saw Mary, that she rose up hastily and went out, followed her, saying, She goeth unto

the grave to weep there. domo et consolabantur eam cum uidissent Mariam quia cito

surrexit et exiit secuti sunt eam dicentes quia uadit ad monu-

11:32 Then when Mary was come where Jesus was, and mentum ut ploret ibi 11:32 Maria ergo cum uenisset ubi erat

saw him, she fell down at his feet, saying unto him, Lord, if thou hadst been here, my brother had not died.

Iesus uidens eum cecidit ad pedes eius et dixit ei Domine si

11:33 When Jesus therefore saw her weeping, and the fuisses hic non esset mortuus frater meus 11:33 Iesus ergo

Jews also weeping which came with her, he groaned in the spirit, and was troubled.

ut uidit eam plorantem et Iudaeos qui uenerant cum ea plo-

11:34 And said, Where have ye laid him? They said unto rantes fremuit spiritu et turbauit se ipsum 11:34 et dixit ubi

him, Lord, come and see.

11:35 Jesus wept. posuistis eum dicunt ei Domine ueni et uide 11:35 et lac-11:36 Then said the Jews, Behold how he loved him! rimatus est Iesus 11:36 dixerunt ergo Iudaei ecce quomodo 11:37 And some of them said, Could not this man, which amabat eum 11:37 quidam autem dixerunt ex ipsis non po-

opened the eyes of the blind, have caused that even this man should not have died?

terat hic qui aperuit oculos caeci facere ut et hic non morere-

11:38 Jesus therefore again groaning in himself cometh to tur 11:38 Iesus ergo rursum fremens in semet ipso uenit ad the grave. It was a cave, and a stone lay upon it.

monumentum erat autem spelunca et lapis superpositus erat

11:39 Jesus said, Take ye away the stone. Martha, the ei 11:39 ait Iesus tollite lapidem dicit ei Martha soror eius

sister of him that was dead, saith unto him, Lord, by this time he stinketh: for he hath been dead four days.

qui mortuus fuerat Domine iam fetet quadriduanus enim est

11:40 Jesus saith unto her, Said I not unto thee, that, if thou 11:40 dicit ei Iesus nonne dixi tibi quoniam si credideris ui-

wouldest believe, thou shouldest see the glory of God?

11:41 Then they took away the stone from the place where debis gloriam Dei 11:41 tulerunt ergo lapidem Iesus autem

the dead was laid. And Jesus lifted up his eyes, and said, Father, I thank thee that thou hast heard me.

eleuatis sursum oculis dixit Pater gratias ago tibi quoniam

11:42 And I knew that thou hearest me always: but be- audisti me 11:42 ego autem sciebam quia semper me audis

cause of the people which stand by I said it, that they may believe that thou hast sent me.

sed propter populum qui circumstat dixi ut credant quia tu

11:43 And when he thus had spoken, he cried with a loud me misisti 11:43 haec cum dixisset uoce magna clamauit

voice, Lazarus, come forth.

11:44 And he that was dead came forth, bound hand and Lazare ueni foras 11:44 et statim prodiit qui fuerat mortuus

foot with graveclothes: and his face was bound about with a napkin. Jesus saith unto them, Loose him, and let him

go. ligatus pedes et manus institis et facies illius sudario erat li-

Euangelium Secundum Ioannem

gata dicit Iesus eis soluite eum et sinite abire 11:45 multi 11:45 Then many of the Jews which came to Mary, and

had seen the things which Jesus did, believed on him.

ergo ex Iudaeis qui uenerant ad Mariam et uiderant quae fecit crediderunt in eum 11:46 quidam autem ex ipsis abierunt 11:46 But some of them went their ways to the Pharisees,

and told them what things Jesus had done.

ad Pharisaeos et dixerunt eis quae fecit Iesus 11:47 college- 11:47 Then gathered the chief priests and the Pharisees a

council, and said, What do we? for this man doeth many miracles.

runt ergo pontifices et Pharisaei concilium et dicebant quid facimus quia hic homo multa signa facit 11:48 si dimitti- 11:48 If we let him thus alone, all men will believe on him:

and the Romans shall come and take away both our place and nation.

mus eum sic omnes credent in eum et uenient Romani et tollent nostrum et locum et gentem 11:49 unus autem ex ipsis 11:49 And one of them, named Caiaphas, being the high

priest that same year, said unto them, Ye know nothing at all,

Caiaphas cum esset pontifex anni illius dixit eis uos nescitis quicquam 11:50 nec cogitatis quia expedit nobis ut unus 11:50 Nor consider that it is expedient for us, that one man

should die for the people, and that the whole nation perish not.

moriatur homo pro populo et non tota gens pereat 11:51 hoc

11:51 And this spake he not of himself: but being high priest that year, he prophesied that Jesus should die for that nation;

autem a semet ipso non dixit sed cum esset pontifex anni illius prophetauit quia Iesus moriturus erat pro gente 11:52 11:52 And not for that nation only, but that also he should

gather together in one the children of God that were scattered abroad.

et non tantum pro gente sed et ut filios Dei qui erant dispersi congregaret in unum 11:53 ab illo ergo die cogitauerunt ut 11:53 Then from that day forth they took counsel together

for to put him to death.

interficerent eum 11:54 Iesus ergo iam non in palam ambu- 11:54 Jesus therefore walked no more openly among the

Jews; but went thence unto a country near to the wilderness, into a city called Ephraim, and there continued with

labat apud Iudaeos sed abiit in regionem iuxta desertum in his disciples. ciuitatem quae dicitur Efrem et ibi morabatur cum discipulis 11:55 proximum autem erat pascha Iudaeorum et ascenderunt 11:55 And the Jews’ passover was nigh at hand: and

many went out of the country up to Jerusalem before the passover, to purify themselves.

multi Hierosolyma de regione ante pascha ut sanctificarent se ipsos 11:56 quaerebant ergo Iesum et conloquebantur ad 11:56 Then sought they for Jesus, and spake among them-

selves, as they stood in the temple, What think ye, that he will not come to the feast?

inuicem in templo stantes quid putatis quia non ueniat ad diem festum 11:57 dederant autem pontifices et Pharisaei 11:57 Now both the chief priests and the Pharisees had

given a commandment, that, if any man knew where he were, he should shew it, that they might take him.

mandatum ut si quis cognouerit ubi sit indicet ut adprehendant eum

12:1 Iesus ergo ante sex dies paschae uenit Bethaniam ubi 12:1 Then Jesus six days before the passover came to

Bethany, where Lazarus was, which had been dead, whom he raised from the dead.

fuerat Lazarus mortuus quem suscitauit Iesus 12:2 fece-

12:2 There they made him a supper; and Martha served: but Lazarus was one of them that sat at the table with him.

runt autem ei cenam ibi et Martha ministrabat Lazarus uero unus erat ex discumbentibus cum eo 12:3 Maria ergo ac- 12:3 Then took Mary a pound of ointment of spikenard,

very costly, and anointed the feet of Jesus, and wiped his feet with her hair: and the house was filled with the odour

cepit libram unguenti nardi pistici pretiosi unxit pedes Iesu of the ointment. et extersit capillis suis pedes eius et domus impleta est ex odore unguenti 12:4 dicit ergo unus ex discipulis eius Iu- 12:4 Then saith one of his disciples, Judas Iscariot, Si-

mon’s son, which should betray him,

das Scariotis qui erat eum traditurus 12:5 quare hoc un- 12:5 Why was not this ointment sold for three hundred

pence, and given to the poor?

Biblia Sacra Vulgata

guentum non ueniit trecentis denariis et datum est egenis

12:6 This he said, not that he cared for the poor; but be- 12:6 dixit autem hoc non quia de egenis pertinebat ad eum

cause he was a thief, and had the bag, and bare what was put therein.

sed quia fur erat et loculos habens ea quae mittebantur por-

12:7 Then said Jesus, Let her alone: against the day of my tabat 12:7 dixit ergo Iesus sine illam ut in die sepulturae

burying hath she kept this.

12:8 For the poor always ye have with you; but me ye have meae seruet illud 12:8 pauperes enim semper habetis uo-

not always. 12:9 Much people of the Jews therefore knew that he was biscum me autem non semper habetis 12:9 cognouit ergo there: and they came not for Jesus’ sake only, but that they might see Lazarus also, whom he had raised from the

dead. turba multa ex Iudaeis quia illic est et uenerunt non prop-

ter Iesum tantum sed ut Lazarum uiderent quem suscitauit a

12:10 But the chief priests consulted that they might put mortuis 12:10 cogitauerunt autem principes sacerdotum ut

Lazarus also to death;

12:11 Because that by reason of him many of the Jews et Lazarum interficerent 12:11 quia multi propter illum abi-

went away, and believed on Jesus.

12:12 On the next day much people that were come to bant ex Iudaeis et credebant in Iesum 12:12 in crastinum

the feast, when they heard that Jesus was coming to Jerusalem,

autem turba multa quae uenerat ad diem festum cum audis-

12:13 Took branches of palm trees, and went forth to meet sent quia uenit Iesus Hierosolyma 12:13 acceperunt ramos

him, and cried, Hosanna: Blessed is the King of Israel that cometh in the name of the Lord.

palmarum et processerunt obuiam ei et clamabant osanna

12:14 And Jesus, when he had found a young ass, sat benedictus qui uenit in nomine Domini rex Israhel 12:14

thereon; as it is written,

et inuenit Iesus asellum et sedit super eum sicut scriptum

12:15 Fear not, daughter of Sion: behold, thy King est 12:15 noli timere filia Sion ecce rex tuus uenit sedens

cometh, sitting on an ass’s colt. 12:16 These things understood not his disciples at the first: super pullum asinae 12:16 haec non cognouerunt discipuli but when Jesus was glorified, then remembered they that these things were written of him, and that they had done

these things unto him. eius primum sed quando glorificatus est Iesus tunc recordati 12:17 The people therefore that was with him when he sunt quia haec erant scripta de eo et haec fecerunt ei 12:17

called Lazarus out of his grave, and raised him from the dead, bare record.

testimonium ergo perhibebat turba quae erat cum eo quando Lazarum uocauit de monumento et suscitauit eum a mor-

12:18 For this cause the people also met him, for that they tuis 12:18 propterea et obuiam uenit ei turba quia audierunt

heard that he had done this miracle.

12:19 The Pharisees therefore said among themselves, eum fecisse hoc signum 12:19 Pharisaei ergo dixerunt ad

Perceive ye how ye prevail nothing? behold, the world is gone after him.

semet ipsos uidetis quia nihil proficimus ecce mundus totus

12:20 And there were certain Greeks among them that post eum abiit 12:20 erant autem gentiles quidam ex his qui

came up to worship at the feast:

12:21 The same came therefore to Philip, which was of ascenderant ut adorarent in die festo 12:21 hii ergo acces-

Bethsaida of Galilee, and desired him, saying, Sir, we would see Jesus.

serunt ad Philippum qui erat a Bethsaida Galilaeae et roga-

12:22 Philip cometh and telleth Andrew: and again An- bant eum dicentes domine uolumus Iesum uidere 12:22 ue-

drew and Philip tell Jesus.

nit Philippus et dicit Andreae Andreas rursum et Philippus

12:23 And Jesus answered them, saying, The hour is dixerunt Iesu 12:23 Iesus autem respondit eis dicens uenit

come, that the Son of man should be glorified. 12:24 Verily, verily, I say unto you, Except a corn of wheat hora ut clarificetur Filius hominis 12:24 amen amen dico fall into the ground and die, it abideth alone: but if it die, it bringeth forth much fruit.

uobis nisi granum frumenti cadens in terram mortuum fuerit

Euangelium Secundum Ioannem

12:25 ipsum solum manet si autem mortuum fuerit multum 12:25 He that loveth his life shall lose it; and he that hateth

his life in this world shall keep it unto life eternal.

fructum adfert qui amat animam suam perdet eam et qui odit animam suam in hoc mundo in uitam aeternam custodit eam 12:26 si quis mihi ministrat me sequatur et ubi sum ego illic 12:26 If any man serve me, let him follow me; and where

I am, there shall also my servant be: if any man serve me, him will my Father honour.

et minister meus erit si quis mihi ministrauerit honorifica-

bit eum Pater meus 12:27 nunc anima mea turbata est et 12:27 Now is my soul troubled; and what shall I say? Father, save me from this hour: but for this cause came I unto this hour.

quid dicam Pater saluifica me ex hora hac sed propterea ueni in horam hanc 12:28 Pater clarifica tuum nomen uenit ergo 12:28 Father, glorify thy name. Then came there a voice

from heaven, saying, I have both glorified it, and will glorify it again.

uox de caelo et clarificaui et iterum clarificabo 12:29 turba

12:29 The people therefore, that stood by, and heard it, said that it thundered: others said, An angel spake to him.

ergo quae stabat et audierat dicebant tonitruum factum esse alii dicebant angelus ei locutus est 12:30 respondit Iesus et 12:30 Jesus answered and said, This voice came not be-

cause of me, but for your sakes.

dixit non propter me uox haec uenit sed propter uos 12:31 12:31 Now is the judgment of this world: now shall the

prince of this world be cast out.

nunc iudicium est mundi nunc princeps huius mundi eicietur foras 12:32 et ego si exaltatus fuero a terra omnia traham 12:32 And I, if I be lifted up from the earth, will draw all

men unto me.

ad me ipsum 12:33 hoc autem dicebat significans qua morte 12:33 This he said, signifying what death he should die. esset moriturus 12:34 respondit ei turba nos audiuimus ex 12:34 The people answered him, We have heard out of the

law that Christ abideth for ever: and how sayest thou, The Son of man must be lifted up? who is this Son of man?

lege quia Christus manet in aeternum et quomodo tu dicis oportet exaltari Filium hominis quis est iste Filius hominis 12:35 dixit ergo eis Iesus adhuc modicum lumen in uobis est 12:35 Then Jesus said unto them, Yet a little while is the

light with you. Walk while ye have the light, lest darkness come upon you: for he that walketh in darkness knoweth

ambulate dum lucem habetis ut non tenebrae uos conprehen- not whither he goeth.

dant et qui ambulat in tenebris nescit quo uadat 12:36 dum 12:36 While ye have light, believe in the light, that ye may be the children of light. These things spake Jesus, and departed, and did hide himself from them.

lucem habetis credite in lucem ut filii lucis sitis haec locutus est Iesus et abiit et abscondit se ab eis 12:37 cum autem 12:37 But though he had done so many miracles before

them, yet they believed not on him:

tanta signa fecisset coram eis non credebant in eum 12:38 ut 12:38 That the saying of Esaias the prophet might be ful-

filled, which he spake, Lord, who hath believed our report? and to whom hath the arm of the Lord been revealed?

sermo Esaiae prophetae impleretur quem dixit Domine quis credidit auditui nostro et brachium Domini cui reuelatum est 12:39 propterea non poterant credere quia iterum dixit Esaias 12:39 Therefore they could not believe, because that Esa-

ias said again,

12:40 excaecauit oculos eorum et indurauit eorum cor ut non 12:40 He hath blinded their eyes, and hardened their heart;

that they should not see with their eyes, nor understand with their heart, and be converted, and I should heal them.

uideant oculis et intellegant corde et conuertantur et sanem eos 12:41 haec dixit Esaias quando uidit gloriam eius et lo- 12:41 These things said Esaias, when he saw his glory, and

spake of him.

cutus est de eo 12:42 uerumtamen et ex principibus multi 12:42 Nevertheless among the chief rulers also many be-

lieved on him; but because of the Pharisees they did not confess him, lest they should be put out of the synagogue:

crediderunt in eum sed propter Pharisaeos non confitebantur ut de synagoga non eicerentur 12:43 dilexerunt enim glo- 12:43 For they loved the praise of men more than the

praise of God.

Biblia Sacra Vulgata

12:44 Jesus cried and said, He that believeth on me, be- riam hominum magis quam gloriam Dei 12:44 Iesus autem lieveth not on me, but on him that sent me.

clamauit et dixit qui credit in me non credit in me sed in

12:45 And he that seeth me seeth him that sent me. eum qui misit me 12:45 et qui uidet me uidet eum qui mi-

12:46 I am come a light into the world, that whosoever sit me 12:46 ego lux in mundum ueni ut omnis qui credit believeth on me should not abide in darkness.

12:47 And if any man hear my words, and believe not, I in me in tenebris non maneat 12:47 et si quis audierit uerba

judge him not: for I came not to judge the world, but to save the world.

mea et non custodierit ego non iudico eum non enim ueni

12:48 He that rejecteth me, and receiveth not my words, ut iudicem mundum sed ut saluificem mundum 12:48 qui hath one that judgeth him: the word that I have spoken, the same shall judge him in the last day.

spernit me et non accipit uerba mea habet qui iudicet eum sermo quem locutus sum ille iudicabit eum in nouissimo die

12:49 For I have not spoken of myself; but the Father 12:49 quia ego ex me ipso non sum locutus sed qui misit me

which sent me, he gave me a commandment, what I should say, and what I should speak.

Pater ipse mihi mandatum dedit quid dicam et quid loquar

12:50 And I know that his commandment is life everlast- 12:50 et scio quia mandatum eius uita aeterna est quae ergo

ing: whatsoever I speak therefore, even as the Father said unto me, so I speak.

ego loquor sicut dixit mihi Pater sic loquor

13:1 Now before the feast of the passover, when Jesus 13:1 ante diem autem festum paschae sciens Iesus quia ue-

knew that his hour was come that he should depart out of this world unto the Father, having loved his own which

were in the world, he loved them unto the end. nit eius hora ut transeat ex hoc mundo ad Patrem cum di-13:2 And supper being ended, the devil having now put lexisset suos qui erant in mundo in finem dilexit eos 13:2

into the heart of Judas Iscariot, Simon’s son, to betray him;

et cena facta cum diabolus iam misisset in corde ut trade-

13:3 Jesus knowing that the Father had given all things ret eum Iudas Simonis Scariotis 13:3 sciens quia omnia de-

into his hands, and that he was come from God, and went to God;

dit ei Pater in manus et quia a Deo exiuit et ad Deum uadit

13:4 He riseth from supper, and laid aside his garments; 13:4 surgit a cena et ponit uestimenta sua et cum accepis-

and took a towel, and girded himself.

13:5 After that he poureth water into a bason, and began to set linteum praecinxit se 13:5 deinde mittit aquam in pel-

wash the disciples’ feet, and to wipe them with the towel wherewith he was girded.

uem et coepit lauare pedes discipulorum et extergere linteo

13:6 Then cometh he to Simon Peter: and Peter saith unto quo erat praecinctus 13:6 uenit ergo ad Simonem Petrum et

him, Lord, dost thou wash my feet?

13:7 Jesus answered and said unto him, What I do thou dicit ei Petrus Domine tu mihi lauas pedes 13:7 respondit

knowest not now; but thou shalt know hereafter.

Iesus et dicit ei quod ego facio tu nescis modo scies autem

13:8 Peter saith unto him, Thou shalt never wash my feet. postea 13:8 dicit ei Petrus non lauabis mihi pedes in aeter-

Jesus answered him, If I wash thee not, thou hast no part with me.

num respondit Iesus ei si non lauero te non habes partem

13:9 Simon Peter saith unto him, Lord, not my feet only, mecum 13:9 dicit ei Simon Petrus Domine non tantum pebut also my hands and my head.

13:10 Jesus saith to him, He that is washed needeth not des meos sed et manus et caput 13:10 dicit ei Iesus qui lotus

save to wash his feet, but is clean every whit: and ye are clean, but not all.

est non indiget ut lauet sed est mundus totus et uos mundi

13:11 For he knew who should betray him; therefore said estis sed non omnes 13:11 sciebat enim quisnam esset qui

he, Ye are not all clean.

13:12 So after he had washed their feet, and had taken traderet eum propterea dixit non estis mundi omnes 13:12

his garments, and was set down again, he said unto them, Know ye what I have done to you?

postquam ergo lauit pedes eorum et accepit uestimenta sua

Euangelium Secundum Ioannem

cum recubuisset iterum dixit eis scitis quid fecerim uobis 13:13 uos uocatis me magister et Domine et bene dicitis sum 13:13 Ye call me Master and Lord: and ye say well; for so

I am.

etenim 13:14 si ergo ego laui uestros pedes Dominus et ma- 13:14 If I then, your Lord and Master, have washed your

feet; ye also ought to wash one another’s feet.

gister et uos debetis alter alterius lauare pedes 13:15 exem- 13:15 For I have given you an example, that ye should do

as I have done to you.

plum enim dedi uobis ut quemadmodum ego feci uobis ita et uos faciatis 13:16 amen amen dico uobis non est seruus 13:16 Verily, verily, I say unto you, The servant is not

greater than his lord; neither he that is sent greater than he that sent him.

maior domino suo neque apostolus maior eo qui misit illum 13:17 si haec scitis beati eritis si feceritis ea 13:18 non de 13:17 If ye know these things, happy are ye if ye do them.

13:18 I speak not of you all: I know whom I have chosen: but that the scripture may be fulfilled, He that eateth bread with me hath lifted up his heel against me.

omnibus uobis dico ego scio quos elegerim sed ut impleatur scriptura qui manducat mecum panem leuauit contra me calcaneum suum 13:19 amodo dico uobis priusquam fiat ut 13:19 Now I tell you before it come, that, when it is come

to pass, ye may believe that I am he.

credatis cum factum fuerit quia ego sum 13:20 amen amen 13:20 Verily, verily, I say unto you, He that receiveth

whomsoever I send receiveth me; and he that receiveth me receiveth him that sent me.

dico uobis qui accipit si quem misero me accipit qui autem

me accipit accipit eum qui me misit 13:21 cum haec dixis- 13:21 When Jesus had thus said, he was troubled in spirit, and testified, and said, Verily, verily, I say unto you, that one of you shall betray me.

set Iesus turbatus est spiritu et protestatus est et dixit amen

amen dico uobis quia unus ex uobis tradet me 13:22 aspi- 13:22 Then the disciples looked one on another, doubting

of whom he spake.

ciebant ergo ad inuicem discipuli haesitantes de quo diceret 13:23 erat ergo recumbens unus ex discipulis eius in sinu Iesu 13:23 Now there was leaning on Jesus’ bosom one of his

disciples, whom Jesus loved.

quem diligebat Iesus 13:24 innuit ergo huic Simon Petrus 13:24 Simon Peter therefore beckoned to him, that he

should ask who it should be of whom he spake.

et dicit ei quis est de quo dicit 13:25 itaque cum recubuisset 13:25 He then lying on Jesus’ breast saith unto him, Lord,

who is it?

ille supra pectus Iesu dicit ei Domine quis est 13:26 res- 13:26 Jesus answered, He it is, to whom I shall give a sop,

when I have dipped it. And when he had dipped the sop, he gave it to Judas Iscariot, the son of Simon.

pondit Iesus ille est cui ego intinctum panem porrexero et cum intinxisset panem dedit Iudae Simonis Scariotis 13:27 13:27 And after the sop Satan entered into him. Then said

Jesus unto him, That thou doest, do quickly.

et post buccellam tunc introiuit in illum Satanas dicit ei Iesus quod facis fac citius 13:28 hoc autem nemo sciuit di- 13:28 Now no man at the table knew for what intent he

spake this unto him.

scumbentium ad quid dixerit ei 13:29 quidam enim puta- 13:29 For some of them thought, because Judas had the

bag, that Jesus had said unto him, Buy those things that we have need of against the feast; or, that he should give

bant quia loculos habebat Iudas quia dicit ei Iesus eme ea something to the poor. quae opus sunt nobis ad diem festum aut egenis ut aliquid daret 13:30 cum ergo accepisset ille buccellam exiuit conti- 13:30 He then having received the sop went immediately

out: and it was night.

nuo erat autem nox 13:31 cum ergo exisset dicit Iesus nunc 13:31 Therefore, when he was gone out, Jesus said, Now

is the Son of man glorified, and God is glorified in him.

clarificatus est Filius hominis et Deus clarificatus est in eo 13:32 si Deus clarificatus est in eo et Deus clarificabit eum in 13:32 If God be glorified in him, God shall also glorify

him in himself, and shall straightway glorify him.

semet ipso et continuo clarificabit eum 13:33 filioli adhuc 13:33 Little children, yet a little while I am with you. Ye

shall seek me: and as I said unto the Jews, Whither I go, ye cannot come; so now I say to you.

Biblia Sacra Vulgata

modicum uobiscum sum quaeretis me et sicut dixi Iudaeis quo ego uado uos non potestis uenire et uobis dico modo

13:34 A new commandment I give unto you, That ye love 13:34 mandatum nouum do uobis ut diligatis inuicem sicut

one another; as I have loved you, that ye also love one another.

dilexi uos ut et uos diligatis inuicem 13:35 in hoc cognos-

13:35 By this shall all men know that ye are my disciples, if ye have love one to another.

cent omnes quia mei discipuli estis si dilectionem habueritis

13:36 Simon Peter said unto him, Lord, whither goest ad inuicem 13:36 dicit ei Simon Petrus Domine quo uadis

thou? Jesus answered him, Whither I go, thou canst not follow me now; but thou shalt follow me afterwards.

respondit Iesus quo ego uado non potes me modo sequi se-

13:37 Peter said unto him, Lord, why cannot I follow thee queris autem postea 13:37 dicit ei Petrus quare non possum

now? I will lay down my life for thy sake. 13:38 Jesus answered him, Wilt thou lay down thy life for sequi te modo animam meam pro te ponam 13:38 respondit my sake? Verily, verily, I say unto thee, The cock shall not crow, till thou hast denied me thrice.

Iesus animam tuam pro me ponis amen amen dico tibi non cantabit gallus donec me ter neges

14:1 Let not your heart be troubled: ye believe in God, 14:1 non turbetur cor uestrum creditis in Deum et in me believe also in me.

14:2 In my Father’s house are many mansions: if it were credite 14:2 in domo Patris mei mansiones multae sunt si

not so, I would have told you. I go to prepare a place for you.

quo minus dixissem uobis quia uado parare uobis locum

14:3 And if I go and prepare a place for you, I will come 14:3 et si abiero et praeparauero uobis locum iterum uenio et

again, and receive you unto myself; that where I am, there ye may be also.

accipiam uos ad me ipsum ut ubi sum ego et uos sitis 14:4 et

14:4 And whither I go ye know, and the way ye know.

14:5 Thomas saith unto him, Lord, we know not whither quo ego uado scitis et uiam scitis 14:5 dicit ei Thomas Do-

thou goest; and how can we know the way?

mine nescimus quo uadis et quomodo possumus uiam scire

14:6 Jesus saith unto him, I am the way, the truth, and the 14:6 dicit ei Iesus ego sum uia et ueritas et uita nemo uenit

life: no man cometh unto the Father, but by me.

14:7 If ye had known me, ye should have known my Father ad Patrem nisi per me 14:7 si cognouissetis me et Patrem

also: and from henceforth ye know him, and have seen him.

meum utique cognouissetis et amodo cognoscitis eum et ui-

14:8 Philip saith unto him, Lord, shew us the Father, and distis eum 14:8 dicit ei Philippus Domine ostende nobis Pa-

it sufficeth us.

14:9 Jesus saith unto him, Have I been so long time with trem et sufficit nobis 14:9 dicit ei Iesus tanto tempore uobi-

you, and yet hast thou not known me, Philip? he that hath seen me hath seen the Father; and how sayest thou then,

Shew us the Father? scum sum et non cognouistis me Philippe qui uidit me uidit 14:10 Believest thou not that I am in the Father, and the et Patrem quomodo tu dicis ostende nobis Patrem 14:10 non

Father in me? the words that I speak unto you I speak not of myself: but the Father that dwelleth in me, he doeth the

works. credis quia ego in Patre et Pater in me est uerba quae ego lo-

quor uobis a me ipso non loquor Pater autem in me manens

14:11 Believe me that I am in the Father, and the Father in ipse facit opera 14:11 non creditis quia ego in Patre et Pa-

me: or else believe me for the very works’ sake.

14:12 Verily, verily, I say unto you, He that believeth on ter in me est 14:12 alioquin propter opera ipsa credite amen

me, the works that I do shall he do also; and greater works than these shall he do; because I go unto my Father.

amen dico uobis qui credit in me opera quae ego facio et ipse faciet et maiora horum faciet quia ego ad Patrem uado

14:13 And whatsoever ye shall ask in my name, that will I 14:13 et quodcumque petieritis in nomine meo hoc faciam

do, that the Father may be glorified in the Son.

14:14 If ye shall ask any thing in my name, I will do it. ut glorificetur Pater in Filio 14:14 si quid petieritis me in

Euangelium Secundum Ioannem

nomine meo hoc faciam 14:15 si diligitis me mandata mea 14:15 If ye love me, keep my commandments. seruate 14:16 et ego rogabo Patrem et alium paracletum da- 14:16 And I will pray the Father, and he shall give you

another Comforter, that he may abide with you for ever;

bit uobis ut maneat uobiscum in aeternum 14:17 Spiritum 14:17 Even the Spirit of truth; whom the world cannot

receive, because it seeth him not, neither knoweth him: but ye know him; for he dwelleth with you, and shall be in

ueritatis quem mundus non potest accipere quia non uidet you. eum nec scit eum uos autem cognoscitis eum quia apud uos

manebit et in uobis erit 14:18 non relinquam uos orfanos 14:18 I will not leave you comfortless: I will come to you. ueniam ad uos 14:19 adhuc modicum et mundus me iam 14:19 Yet a little while, and the world seeth me no more;

but ye see me: because I live, ye shall live also.

non uidet uos autem uidetis me quia ego uiuo et uos uiuetis 14:20 in illo die uos cognoscetis quia ego sum in Patre meo 14:20 At that day ye shall know that I am in my Father,

and ye in me, and I in you.

et uos in me et ego in uobis 14:21 qui habet mandata mea 14:21 He that hath my commandments, and keepeth them, he it is that loveth me: and he that loveth me shall be loved of my Father, and I will love him, and will manifest myself

et seruat ea ille est qui diligit me qui autem diligit me dilige- to him. tur a Patre meo et ego diligam eum et manifestabo ei me ipsum 14:22 dicit ei Iudas non ille Scariotis Domine quid fac- 14:22 Judas saith unto him, not Iscariot, Lord, how is it

that thou wilt manifest thyself unto us, and not unto the world?

tum est quia nobis manifestaturus es te ipsum et non mundo 14:23 respondit Iesus et dixit ei si quis diligit me sermonem 14:23 Jesus answered and said unto him, If a man love me,

he will keep my words: and my Father will love him, and we will come unto him, and make our abode with him.

meum seruabit et Pater meus diliget eum et ad eum ueniemus et mansiones apud eum faciemus 14:24 qui non diligit me 14:24 He that loveth me not keepeth not my sayings: and

the word which ye hear is not mine, but the Father’s which sent me.

sermones meos non seruat et sermonem quem audistis non est meus sed eius qui misit me Patris 14:25 haec locutus 14:25 These things have I spoken unto you, being yet

present with you.

sum uobis apud uos manens 14:26 paracletus autem Spiri- 14:26 But the Comforter, which is the Holy Ghost, whom

the Father will send in my name, he shall teach you all things, and bring all things to your remembrance, whatso-

tus Sanctus quem mittet Pater in nomine meo ille uos doce- ever I have said unto you. bit omnia et suggeret uobis omnia quaecumque dixero uobis 14:27 pacem relinquo uobis pacem meam do uobis non quo- 14:27 Peace I leave with you, my peace I give unto you:

not as the world giveth, give I unto you. Let not your heart be troubled, neither let it be afraid.

modo mundus dat ego do uobis non turbetur cor uestrum neque formidet 14:28 audistis quia ego dixi uobis uado et 14:28 Ye have heard how I said unto you, I go away, and

come again unto you. If ye loved me, ye would rejoice, because I said, I go unto the Father: for my Father is greater

uenio ad uos si diligeretis me gauderetis utique quia uado than I. ad Patrem quia Pater maior me est 14:29 et nunc dixi uobis 14:29 And now I have told you before it come to pass, that,

when it is come to pass, ye might believe.

priusquam fiat ut cum factum fuerit credatis 14:30 iam non 14:30 Hereafter I will not talk much with you: for the

prince of this world cometh, and hath nothing in me.

multa loquar uobiscum uenit enim princeps mundi huius et

in me non habet quicquam 14:31 sed ut cognoscat mundus 14:31 But that the world may know that I love the Father; and as the Father gave me commandment, even so I do. Arise, let us go hence.

quia diligo Patrem et sicut mandatum dedit mihi Pater sic facio surgite eamus hinc

15:1 ego sum uitis uera et Pater meus agricola est 15:2 15:1 I am the true vine, and my Father is the husbandman. 15:2 Every branch in me that beareth not fruit he taketh away: and every branch that beareth fruit, he purgeth it, that it may bring forth more fruit.

Biblia Sacra Vulgata

omnem palmitem in me non ferentem fructum tollet eum et omnem qui fert fructum purgabit eum ut fructum plus adfe-

15:3 Now ye are clean through the word which I have spo- rat 15:3 iam uos mundi estis propter sermonem quem loken unto you. 15:4 Abide in me, and I in you. As the branch cannot bear cutus sum uobis 15:4 manete in me et ego in uobis sicut fruit of itself, except it abide in the vine; no more can ye, except ye abide in me.

palmes non potest ferre fructum a semet ipso nisi manserit

15:5 I am the vine, ye are the branches: He that abideth in in uite sic nec uos nisi in me manseritis 15:5 ego sum uitis

me, and I in him, the same bringeth forth much fruit: for without me ye can do nothing.

uos palmites qui manet in me et ego in eo hic fert fructum

15:6 If a man abide not in me, he is cast forth as a branch, multum quia sine me nihil potestis facere 15:6 si quis in

and is withered; and men gather them, and cast them into the fire, and they are burned.

me non manserit mittetur foras sicut palmes et aruit et colli-

15:7 If ye abide in me, and my words abide in you, ye shall gent eos et in ignem mittunt et ardent 15:7 si manseritis in

ask what ye will, and it shall be done unto you.

me et uerba mea in uobis manserint quodcumque uolueritis

15:8 Herein is my Father glorified, that ye bear much fruit; petetis et fiet uobis 15:8 in hoc clarificatus est Pater meus so shall ye be my disciples.

ut fructum plurimum adferatis et efficiamini mei discipuli

15:9 As the Father hath loved me, so have I loved you: 15:9 sicut dilexit me Pater et ego dilexi uos manete in di-

continue ye in my love.

15:10 If ye keep my commandments, ye shall abide in my lectione mea 15:10 si praecepta mea seruaueritis manebitis

love; even as I have kept my Father’s commandments, and abide in his love.

in dilectione mea sicut et ego Patris mei praecepta seruaui

15:11 These things have I spoken unto you, that my joy et maneo in eius dilectione 15:11 haec locutus sum uobis

might remain in you, and that your joy might be full.

ut gaudium meum in uobis sit et gaudium uestrum imple-

15:12 This is my commandment, That ye love one another, atur 15:12 hoc est praeceptum meum ut diligatis inuicem

as I have loved you.

15:13 Greater love hath no man than this, that a man lay sicut dilexi uos 15:13 maiorem hac dilectionem nemo ha-

down his life for his friends.

15:14 Ye are my friends, if ye do whatsoever I command bet ut animam suam quis ponat pro amicis suis 15:14 uos

you.

15:15 Henceforth I call you not servants; for the servant amici mei estis si feceritis quae ego praecipio uobis 15:15

knoweth not what his lord doeth: but I have called you friends; for all things that I have heard of my Father I have

made known unto you. iam non dico uos seruos quia seruus nescit quid facit do-

minus eius uos autem dixi amicos quia omnia quaecumque

15:16 Ye have not chosen me, but I have chosen you, and audiui a Patre meo nota feci uobis 15:16 non uos me elegis-

ordained you, that ye should go and bring forth fruit, and that your fruit should remain: that whatsoever ye shall ask

of the Father in my name, he may give it you. tis sed ego elegi uos et posui uos ut eatis et fructum adferatis

et fructus uester maneat ut quodcumque petieritis Patrem in

15:17 These things I command you, that ye love one an- nomine meo det uobis 15:17 haec mando uobis ut diligatis

other.

15:18 If the world hate you, ye know that it hated me be- inuicem 15:18 si mundus uos odit scitote quia me priorem

fore it hated you.

15:19 If ye were of the world, the world would love his uobis odio habuit 15:19 si de mundo fuissetis mundus quod

own: but because ye are not of the world, but I have chosen you out of the world, therefore the world hateth you.

suum erat diligeret quia uero de mundo non estis sed ego

15:20 Remember the word that I said unto you, The ser- elegi uos de mundo propterea odit uos mundus 15:20 me-

vant is not greater than his lord. If they have persecuted me, they will also persecute you; if they have kept my say-

ing, they will keep yours also. mentote sermonis mei quem ego dixi uobis non est seruus

Euangelium Secundum Ioannem

maior domino suo si me persecuti sunt et uos persequentur si sermonem meum seruauerunt et uestrum seruabunt 15:21 15:21 But all these things will they do unto you for my

name’s sake, because they know not him that sent me.

sed haec omnia facient uobis propter nomen meum quia nesciunt eum qui misit me 15:22 si non uenissem et locutus 15:22 If I had not come and spoken unto them, they had

not had sin: but now they have no cloak for their sin.

fuissem eis peccatum non haberent nunc autem excusationem non habent de peccato suo 15:23 qui me odit et Patrem 15:23 He that hateth me hateth my Father also. meum odit 15:24 si opera non fecissem in eis quae nemo 15:24 If I had not done among them the works which none

other man did, they had not had sin: but now have they both seen and hated both me and my Father.

alius fecit peccatum non haberent nunc autem et uiderunt

et oderunt et me et Patrem meum 15:25 sed ut impleatur 15:25 But this cometh to pass, that the word might be fulfilled that is written in their law, They hated me without a cause.

sermo qui in lege eorum scriptus est quia odio me habuerunt gratis 15:26 cum autem uenerit paracletus quem ego 15:26 But when the Comforter is come, whom I will send

unto you from the Father, even the Spirit of truth, which proceedeth from the Father, he shall testify of me:

mittam uobis a Patre Spiritum ueritatis qui a Patre procedit ille testimonium perhibebit de me 15:27 et uos testimonium 15:27 And ye also shall bear witness, because ye have

been with me from the beginning.

perhibetis quia ab initio mecum estis

16:1 haec locutus sum uobis ut non scandalizemini 16:2 16:1 These things have I spoken unto you, that ye should

not be offended.

16:2 They shall put you out of the synagogues: yea, the time cometh, that whosoever killeth you will think that he doeth God service.

absque synagogis facient uos sed uenit hora ut omnis qui interficit uos arbitretur obsequium se praestare Deo 16:3 et 16:3 And these things will they do unto you, because they

have not known the Father, nor me.

haec facient quia non nouerunt Patrem neque me 16:4 sed 16:4 But these things have I told you, that when the time

shall come, ye may remember that I told you of them. And these things I said not unto you at the beginning, because

haec locutus sum uobis ut cum uenerit hora eorum reminis- I was with you. camini quia ego dixi uobis 16:5 haec autem uobis ab initio 16:5 But now I go my way to him that sent me; and none

of you asketh me, Whither goest thou?

non dixi quia uobiscum eram at nunc uado ad eum qui me

misit et nemo ex uobis interrogat me quo uadis 16:6 sed 16:6 But because I have said these things unto you, sorrow

hath filled your heart.

quia haec locutus sum uobis tristitia impleuit cor uestrum 16:7 sed ego ueritatem dico uobis expedit uobis ut ego uadam 16:7 Nevertheless I tell you the truth; It is expedient for

you that I go away: for if I go not away, the Comforter will not come unto you; but if I depart, I will send him

si enim non abiero paracletus non ueniet ad uos si autem unto you. abiero mittam eum ad uos 16:8 et cum uenerit ille arguet 16:8 And when he is come, he will reprove the world of

sin, and of righteousness, and of judgment:

mundum de peccato et de iustitia et de iudicio 16:9 de pec- 16:9 Of sin, because they believe not on me; cato quidem quia non credunt in me 16:10 de iustitia uero 16:10 Of righteousness, because I go to my Father, and ye

see me no more;

quia ad Patrem uado et iam non uidebitis me 16:11 de iudi- 16:11 Of judgment, because the prince of this world is

judged.

cio autem quia princeps mundi huius iudicatus est 16:12 ad- 16:12 I have yet many things to say unto you, but ye can-

not bear them now.

huc multa habeo uobis dicere sed non potestis portare modo 16:13 cum autem uenerit ille Spiritus ueritatis docebit uos 16:13 Howbeit when he, the Spirit of truth, is come, he

will guide you into all truth: for he shall not speak of himself; but whatsoever he shall hear, that shall he speak: and

in omnem ueritatem non enim loquetur a semet ipso sed he will shew you things to come.

Biblia Sacra Vulgata

quaecumque audiet loquetur et quae uentura sunt adnuntiabit

16:14 He shall glorify me: for he shall receive of mine, uobis 16:14 ille me clarificabit quia de meo accipiet et ad-

and shall shew it unto you. 16:15 All things that the Father hath are mine: therefore nuntiabit uobis 16:15 omnia quaecumque habet Pater mea said I, that he shall take of mine, and shall shew it unto you.

sunt propterea dixi quia de meo accipit et adnuntiabit uobis

16:16 A little while, and ye shall not see me: and again, a 16:16 modicum et iam non uidebitis me et iterum modicum little while, and ye shall see me, because I go to the Father.

16:17 Then said some of his disciples among themselves, et uidebitis me quia uado ad Patrem 16:17 dixerunt ergo ex

What is this that he saith unto us, A little while, and ye shall not see me: and again, a little while, and ye shall see

me: and, Because I go to the Father? discipulis eius ad inuicem quid est hoc quod dicit nobis mo-

dicum et non uidebitis me et iterum modicum et uidebitis

16:18 They said therefore, What is this that he saith, A me et quia uado ad Patrem 16:18 dicebant ergo quid est hoc

little while? we cannot tell what he saith.

16:19 Now Jesus knew that they were desirous to ask him, quod dicit modicum nescimus quid loquitur 16:19 cognouit

and said unto them, Do ye enquire among yourselves of that I said, A little while, and ye shall not see me: and

again, a little while, and ye shall see me? autem Iesus quia uolebant eum interrogare et dixit eis de hoc

quaeritis inter uos quia dixi modicum et non uidebitis me

16:20 Verily, verily, I say unto you, That ye shall weep et iterum modicum et uidebitis me 16:20 amen amen dico and lament, but the world shall rejoice: and ye shall be sorrowful, but your sorrow shall be turned into joy.

uobis quia plorabitis et flebitis uos mundus autem gaudebit uos autem contristabimini sed tristitia uestra uertetur in gau-

16:21 A woman when she is in travail hath sorrow, because dium 16:21 mulier cum parit tristitiam habet quia uenit hora

her hour is come: but as soon as she is delivered of the child, she remembereth no more the anguish, for joy that

a man is born into the world. eius cum autem pepererit puerum iam non meminit pressu-16:22 And ye now therefore have sorrow: but I will see rae propter gaudium quia natus est homo in mundum 16:22

you again, and your heart shall rejoice, and your joy no man taketh from you.

et uos igitur nunc quidem tristitiam habetis iterum autem uidebo uos et gaudebit cor uestrum et gaudium uestrum nemo

16:23 And in that day ye shall ask me nothing. Verily, tollit a uobis 16:23 et in illo die me non rogabitis quicquam

verily, I say unto you, Whatsoever ye shall ask the Father in my name, he will give it you.

amen amen dico uobis si quid petieritis Patrem in nomine

16:24 Hitherto have ye asked nothing in my name: ask, meo dabit uobis 16:24 usque modo non petistis quicquam

and ye shall receive, that your joy may be full.

in nomine meo petite et accipietis ut gaudium uestrum sit

16:25 These things have I spoken unto you in proverbs: plenum 16:25 haec in prouerbiis locutus sum uobis uenit

but the time cometh, when I shall no more speak unto you in proverbs, but I shall shew you plainly of the Father.

hora cum iam non in prouerbiis loquar uobis sed palam de

16:26 At that day ye shall ask in my name: and I say not Patre adnuntiabo uobis 16:26 illo die in nomine meo pete-

unto you, that I will pray the Father for you:

16:27 For the Father himself loveth you, because ye have tis et non dico uobis quia ego rogabo Patrem de uobis 16:27

loved me, and have believed that I came out from God.

ipse enim Pater amat uos quia uos me amastis et credidistis

16:28 I came forth from the Father, and am come into the quia ego a Deo exiui 16:28 exiui a Patre et ueni in mundum

world: again, I leave the world, and go to the Father. 16:29 His disciples said unto him, Lo, now speakest thou iterum relinquo mundum et uado ad Patrem 16:29 dicunt ei plainly, and speakest no proverb.

discipuli eius ecce nunc palam loqueris et prouerbium nul-

16:30 Now are we sure that thou knowest all things, and lum dicis 16:30 nunc scimus quia scis omnia et non opus needest not that any man should ask thee: by this we believe that thou camest forth from God.

Euangelium Secundum Ioannem

est tibi ut quis te interroget in hoc credimus quia a Deo existi 16:31 respondit eis Iesus modo creditis 16:32 ecce ue- 16:31 Jesus answered them, Do ye now believe?

16:32 Behold, the hour cometh, yea, is now come, that ye shall be scattered, every man to his own, and shall leave me alone: and yet I am not alone, because the Father is with me.

nit hora et iam uenit ut dispergamini unusquisque in propria et me solum relinquatis et non sum solus quia Pater mecum est 16:33 haec locutus sum uobis ut in me pacem habeatis 16:33 These things I have spoken unto you, that in me ye

might have peace. In the world ye shall have tribulation: but be of good cheer; I have overcome the world.

in mundo pressuram habetis sed confidite ego uici mundum

17:1 haec locutus est Iesus et subleuatis oculis in caelum 17:1 These words spake Jesus, and lifted up his eyes to

heaven, and said, Father, the hour is come; glorify thy Son, that thy Son also may glorify thee:

dixit Pater uenit hora clarifica Filium tuum ut Filius tuus clarificet te 17:2 sicut dedisti ei potestatem omnis carnis ut 17:2 As thou hast given him power over all flesh, that he

should give eternal life to as many as thou hast given him.

omne quod dedisti ei det eis uitam aeternam 17:3 haec est 17:3 And this is life eternal, that they might know thee the

only true God, and Jesus Christ, whom thou hast sent.

autem uita aeterna ut cognoscant te solum uerum Deum et quem misisti Iesum Christum 17:4 ego te clarificaui super 17:4 I have glorified thee on the earth: I have finished the

work which thou gavest me to do.

terram opus consummaui quod dedisti mihi ut faciam 17:5 17:5 And now, O Father, glorify thou me with thine own

self with the glory which I had with thee before the world was.

et nunc clarifica me tu Pater apud temet ipsum claritatem quam habui priusquam mundus esset apud te 17:6 manife- 17:6 I have manifested thy name unto the men which thou

gavest me out of the world: thine they were, and thou gavest them me; and they have kept thy word.

staui nomen tuum hominibus quos dedisti mihi de mundo tui erant et mihi eos dedisti et sermonem tuum seruauerunt 17:7 nunc cognouerunt quia omnia quae dedisti mihi abs te 17:7 Now they have known that all things whatsoever thou

hast given me are of thee.

sunt 17:8 quia uerba quae dedisti mihi dedi eis et ipsi ac- 17:8 For I have given unto them the words which thou

gavest me; and they have received them, and have known surely that I came out from thee, and they have believed

ceperunt et cognouerunt uere quia a te exiui et crediderunt that thou didst send me.

quia tu me misisti 17:9 ego pro eis rogo non pro mundo 17:9 I pray for them: I pray not for the world, but for them

which thou hast given me; for they are thine.

rogo sed pro his quos dedisti mihi quia tui sunt 17:10 et 17:10 And all mine are thine, and thine are mine; and I am

glorified in them.

mea omnia tua sunt et tua mea sunt et clarificatus sum in eis 17:11 et iam non sum in mundo et hii in mundo sunt et ego 17:11 And now I am no more in the world, but these are in

the world, and I come to thee. Holy Father, keep through thine own name those whom thou hast given me, that they

ad te uenio Pater sancte serua eos in nomine tuo quos de- may be one, as we are. disti mihi ut sint unum sicut et nos 17:12 cum essem cum 17:12 While I was with them in the world, I kept them in

thy name: those that thou gavest me I have kept, and none of them is lost, but the son of perdition; that the scripture

eis ego seruabam eos in nomine tuo quos dedisti mihi custo- might be fulfilled. diui et nemo ex his periuit nisi filius perditionis ut scriptura impleatur 17:13 nunc autem ad te uenio et haec loquor in 17:13 And now come I to thee; and these things I speak in

the world, that they might have my joy fulfilled in themselves.

mundo ut habeant gaudium meum impletum in semet ipsis 17:14 ego dedi eis sermonem tuum et mundus odio eos ha- 17:14 I have given them thy word; and the world hath

hated them, because they are not of the world, even as I am not of the world.

buit quia non sunt de mundo sicut et ego non sum de mundo 17:15 non rogo ut tollas eos de mundo sed ut serues eos ex 17:15 I pray not that thou shouldest take them out of the

world, but that thou shouldest keep them from the evil.

Biblia Sacra Vulgata

17:16 They are not of the world, even as I am not of the malo 17:16 de mundo non sunt sicut et ego non sum de world. 17:17 Sanctify them through thy truth: thy word is truth. mundo 17:17 sanctifica eos in ueritate sermo tuus ueritas 17:18 As thou hast sent me into the world, even so have I est 17:18 sicut me misisti in mundum et ego misi eos in also sent them into the world. 17:19 And for their sakes I sanctify myself, that they also mundum 17:19 et pro eis ego sanctifico me ipsum ut sint might be sanctified through the truth.

17:20 Neither pray I for these alone, but for them also et ipsi sanctificati in ueritate 17:20 non pro his autem rogo

which shall believe on me through their word;

tantum sed et pro eis qui credituri sunt per uerbum eorum in

17:21 That they all may be one; as thou, Father, art in me, me 17:21 ut omnes unum sint sicut tu Pater in me et ego in and I in thee, that they also may be one in us: that the world may believe that thou hast sent me.

te ut et ipsi in nobis unum sint ut mundus credat quia tu me

17:22 And the glory which thou gavest me I have given misisti 17:22 et ego claritatem quam dedisti mihi dedi eis

them; that they may be one, even as we are one:

17:23 I in them, and thou in me, that they may be made ut sint unum sicut nos unum sumus 17:23 ego in eis et tu in

perfect in one; and that the world may know that thou hast sent me, and hast loved them, as thou hast loved me.

me ut sint consummati in unum et cognoscat mundus quia

17:24 Father, I will that they also, whom thou hast given tu me misisti et dilexisti eos sicut me dilexisti 17:24 Pater

me, be with me where I am; that they may behold my glory, which thou hast given me: for thou lovedst me be-

fore the foundation of the world. quos dedisti mihi uolo ut ubi ego sum et illi sint mecum ut ui-

deant claritatem meam quam dedisti mihi quia dilexisti me

17:25 O righteous Father, the world hath not known thee: ante constitutionem mundi 17:25 Pater iuste et mundus te

but I have known thee, and these have known that thou hast sent me.

non cognouit ego autem te cognoui et hii cognouerunt quia

17:26 And I have declared unto them thy name, and will tu me misisti 17:26 et notum feci eis nomen tuum et notum

declare it: that the love wherewith thou hast loved me may be in them, and I in them.

faciam ut dilectio qua dilexisti me in ipsis sit et ego in ipsis

18:1 When Jesus had spoken these words, he went forth 18:1 haec cum dixisset Iesus egressus est cum discipulis

with his disciples over the brook Cedron, where was a garden, into the which he entered, and his disciples.

suis trans torrentem Cedron ubi erat hortus in quem intro-

18:2 And Judas also, which betrayed him, knew the place: iuit ipse et discipuli eius 18:2 sciebat autem et Iudas qui

for Jesus ofttimes resorted thither with his disciples.

tradebat eum ipsum locum quia frequenter Iesus conuene-

18:3 Judas then, having received a band of men and offi- rat illuc cum discipulis suis 18:3 Iudas ergo cum accepisset

cers from the chief priests and Pharisees, cometh thither with lanterns and torches and weapons.

cohortem et a pontificibus et Pharisaeis ministros uenit illuc

18:4 Jesus therefore, knowing all things that should come cum lanternis et facibus et armis 18:4 Iesus itaque sciens upon him, went forth, and said unto them, Whom seek ye?

omnia quae uentura erant super eum processit et dicit eis

18:5 They answered him, Jesus of Nazareth. Jesus saith quem quaeritis 18:5 responderunt ei Iesum Nazarenum di-

unto them, I am he. And Judas also, which betrayed him, stood with them.

cit eis Iesus ego sum stabat autem et Iudas qui tradebat eum

18:6 As soon then as he had said unto them, I am he, they cum ipsis 18:6 ut ergo dixit eis ego sum abierunt retrorsum

went backward, and fell to the ground.

18:7 Then asked he them again, Whom seek ye? And they et ceciderunt in terram 18:7 iterum ergo eos interrogauit

said, Jesus of Nazareth. 18:8 Jesus answered, I have told you that I am he: if there- quem quaeritis illi autem dixerunt Iesum Nazarenum 18:8 fore ye seek me, let these go their way:

respondit Iesus dixi uobis quia ego sum si ergo me quaeritis

18:9 That the saying might be fulfilled, which he spake, sinite hos abire 18:9 ut impleretur sermo quem dixit quia

Of them which thou gavest me have I lost none.

Euangelium Secundum Ioannem

quos dedisti mihi non perdidi ex ipsis quemquam 18:10 18:10 Then Simon Peter having a sword drew it, and smote the high priest’s servant, and cut off his right ear. The servant’s name was Malchus.

Simon ergo Petrus habens gladium eduxit eum et percussit pontificis seruum et abscidit eius auriculam dextram erat autem nomen seruo Malchus 18:11 dixit ergo Iesus Petro 18:11 Then said Jesus unto Peter, Put up thy sword into

the sheath: the cup which my Father hath given me, shall I not drink it?

mitte gladium in uaginam calicem quem dedit mihi Pater non bibam illum 18:12 cohors ergo et tribunus et ministri 18:12 Then the band and the captain and officers of the

Jews took Jesus, and bound him,

Iudaeorum conprehenderunt Iesum et ligauerunt eum 18:13 18:13 And led him away to Annas first; for he was father

in law to Caiaphas, which was the high priest that same year.

et adduxerunt eum ad Annam primum erat enim socer Caiaphae qui erat pontifex anni illius 18:14 erat autem Caiaphas 18:14 Now Caiaphas was he, which gave counsel to the

Jews, that it was expedient that one man should die for the people.

qui consilium dederat Iudaeis quia expedit unum hominem

mori pro populo 18:15 sequebatur autem Iesum Simon Pe- 18:15 And Simon Peter followed Jesus, and so did another disciple: that disciple was known unto the high priest, and went in with Jesus into the palace of the high priest.

trus et alius discipulus discipulus autem ille erat notus pontifici et introiuit cum Iesu in atrium pontificis 18:16 Pet- 18:16 But Peter stood at the door without. Then went out

that other disciple, which was known unto the high priest, and spake unto her that kept the door, and brought in Peter.

rus autem stabat ad ostium foris exiuit ergo discipulus alius qui erat notus pontifici et dixit ostiariae et introduxit Petrum 18:17 dicit ergo Petro ancilla ostiaria numquid et tu ex disci- 18:17 Then saith the damsel that kept the door unto Peter,

Art not thou also one of this man’s disciples? He saith, I am not.

pulis es hominis istius dicit ille non sum 18:18 stabant au-

18:18 And the servants and officers stood there, who had made a fire of coals; for it was cold: and they warmed themselves: and Peter stood with them, and warmed himself.

tem serui et ministri ad prunas quia frigus erat et calefiebant

erat autem cum eis et Petrus stans et calefaciens se 18:19 18:19 The high priest then asked Jesus of his disciples, and

of his doctrine.

pontifex ergo interrogauit Iesum de discipulis suis et de doctrina eius 18:20 respondit ei Iesus ego palam locutus sum 18:20 Jesus answered him, I spake openly to the world; I

ever taught in the synagogue, and in the temple, whither the Jews always resort; and in secret have I said nothing.

mundo ego semper docui in synagoga et in templo quo omnes Iudaei conueniunt et in occulto locutus sum nihil 18:21 18:21 Why askest thou me? ask them which heard me,

what I have said unto them: behold, they know what I said.

quid me interrogas interroga eos qui audierunt quid locutus sum ipsis ecce hii sciunt quae dixerim ego 18:22 haec au- 18:22 And when he had thus spoken, one of the officers

which stood by struck Jesus with the palm of his hand, saying, Answerest thou the high priest so?

tem cum dixisset unus adsistens ministrorum dedit alapam Iesu dicens sic respondes pontifici 18:23 respondit ei Iesus 18:23 Jesus answered him, If I have spoken evil, bear wit-

ness of the evil: but if well, why smitest thou me?

si male locutus sum testimonium perhibe de malo si autem bene quid me caedis 18:24 et misit eum Annas ligatum ad 18:24 Now Annas had sent him bound unto Caiaphas the

high priest.

Caiaphan pontificem 18:25 erat autem Simon Petrus stans 18:25 And Simon Peter stood and warmed himself. They

said therefore unto him, Art not thou also one of his disciples? He denied it, and said, I am not.

et calefaciens se dixerunt ergo ei numquid et tu ex discipu-

lis eius es negauit ille et dixit non sum 18:26 dicit unus ex 18:26 One of the servants of the high priest, being his kinsman whose ear Peter cut off, saith, Did not I see thee in the garden with him?

seruis pontificis cognatus eius cuius abscidit Petrus auricu-

Biblia Sacra Vulgata

18:27 Peter then denied again: and immediately the cock lam nonne ego te uidi in horto cum illo 18:27 iterum ergo crew.

18:28 Then led they Jesus from Caiaphas unto the hall of negauit Petrus et statim gallus cantauit 18:28 adducunt ergo

judgment: and it was early; and they themselves went not into the judgment hall, lest they should be defiled; but that

they might eat the passover. Iesum a Caiapha in praetorium erat autem mane et ipsi non

introierunt in praetorium ut non contaminarentur sed man-

18:29 Pilate then went out unto them, and said, What ac- ducarent pascha 18:29 exiuit ergo Pilatus ad eos foras et

cusation bring ye against this man?

dixit quam accusationem adfertis aduersus hominem hunc

18:30 They answered and said unto him, If he were not a 18:30 responderunt et dixerunt ei si non esset hic malefac-

malefactor, we would not have delivered him up unto thee.

18:31 Then said Pilate unto them, Take ye him, and judge tor non tibi tradidissemus eum 18:31 dixit ergo eis Pilatus

him according to your law. The Jews therefore said unto him, It is not lawful for us to put any man to death:

accipite eum uos et secundum legem uestram iudicate eum dixerunt ergo ei Iudaei nobis non licet interficere quemquam

18:32 That the saying of Jesus might be fulfilled, which he 18:32 ut sermo Iesu impleretur quem dixit significans qua es-

spake, signifying what death he should die.

18:33 Then Pilate entered into the judgment hall again, set morte moriturus 18:33 introiuit ergo iterum in praeto-

and called Jesus, and said unto him, Art thou the King of the Jews?

rium Pilatus et uocauit Iesum et dixit ei tu es rex Iudaeo-

18:34 Jesus answered him, Sayest thou this thing of thy- rum 18:34 et respondit Iesus a temet ipso hoc dicis an alii

self, or did others tell it thee of me?

18:35 Pilate answered, Am I a Jew? Thine own nation and tibi dixerunt de me 18:35 respondit Pilatus numquid ego Iu-

the chief priests have delivered thee unto me: what hast thou done?

daeus sum gens tua et pontifices tradiderunt te mihi quid fe-

18:36 Jesus answered, My kingdom is not of this world: if cisti 18:36 respondit Iesus regnum meum non est de mundo my kingdom were of this world, then would my servants fight, that I should not be delivered to the Jews: but now is

my kingdom not from hence. hoc si ex hoc mundo esset regnum meum ministri mei decer-

tarent ut non traderer Iudaeis nunc autem meum regnum non

18:37 Pilate therefore said unto him, Art thou a king then? est hinc 18:37 dixit itaque ei Pilatus ergo rex es tu respondit

Jesus answered, Thou sayest that I am a king. To this end was I born, and for this cause came I into the world, that I should bear witness unto the truth. Every one that is of the truth heareth my voice.

Iesus tu dicis quia rex sum ego ego in hoc natus sum et ad hoc ueni in mundum ut testimonium perhibeam ueritati om-

18:38 Pilate saith unto him, What is truth? And when he nis qui est ex ueritate audit meam uocem 18:38 dicit ei Pi-

had said this, he went out again unto the Jews, and saith unto them, I find in him no fault at all.

latus quid est ueritas et cum hoc dixisset iterum exiuit ad Iu-

18:39 But ye have a custom, that I should release unto you daeos et dicit eis ego nullam inuenio in eo causam 18:39 est

one at the passover: will ye therefore that I release unto you the King of the Jews?

autem consuetudo uobis ut unum dimittam uobis in pascha

18:40 Then cried they all again, saying, Not this man, but uultis ergo dimittam uobis regem Iudaeorum 18:40 clama-Barabbas. Now Barabbas was a robber.

uerunt rursum omnes dicentes non hunc sed Barabban erat autem Barabbas latro

19:1 Then Pilate therefore took Jesus, and scourged him. 19:1 tunc ergo adprehendit Pilatus Iesum et flagellauit

19:2 And the soldiers platted a crown of thorns, and put it 19:2 et milites plectentes coronam de spinis inposuerunt ca-

on his head, and they put on him a purple robe,

19:3 And said, Hail, King of the Jews! and they smote him piti eius et ueste purpurea circumdederunt eum 19:3 et ue-

with their hands.

niebant ad eum et dicebant haue rex Iudaeorum et dabant ei

Euangelium Secundum Ioannem

alapas 19:4 exiit iterum Pilatus foras et dicit eis ecce ad- 19:4 Pilate therefore went forth again, and saith unto them, Behold, I bring him forth to you, that ye may know that I find no fault in him.

duco uobis eum foras 19:5 ut cognoscatis quia in eo nul-

19:5 Then came Jesus forth, wearing the crown of thorns, and the purple robe. And Pilate saith unto them, Behold the man!

lam causam inuenio et purpureum uestimentum et dicit eis

ecce homo 19:6 cum ergo uidissent eum pontifices et mi- 19:6 When the chief priests therefore and officers saw him, they cried out, saying, Crucify him, crucify him. Pilate saith unto them, Take ye him, and crucify him: for I find

nistri clamabant dicentes crucifige crucifige dicit eis Pila- no fault in him. tus accipite eum uos et crucifigite ego enim non inuenio in eo causam 19:7 responderunt ei Iudaei nos legem habe- 19:7 The Jews answered him, We have a law, and by our

law he ought to die, because he made himself the Son of God.

mus et secundum legem debet mori quia Filium Dei se fecit 19:8 cum ergo audisset Pilatus hunc sermonem magis timuit 19:8 When Pilate therefore heard that saying, he was the

more afraid;

19:9 et ingressus est praetorium iterum et dicit ad Iesum unde 19:9 And went again into the judgment hall, and saith unto

Jesus, Whence art thou? But Jesus gave him no answer.

es tu Iesus autem responsum non dedit ei 19:10 dicit ergo ei 19:10 Then saith Pilate unto him, Speakest thou not unto

me? knowest thou not that I have power to crucify thee, and have power to release thee?

Pilatus mihi non loqueris nescis quia potestatem habeo crucifigere te et potestatem habeo dimittere te 19:11 respondit 19:11 Jesus answered, Thou couldest have no power at all

against me, except it were given thee from above: therefore he that delivered me unto thee hath the greater sin.

Iesus non haberes potestatem aduersum me ullam nisi tibi esset datum desuper propterea qui tradidit me tibi maius peccatum habet 19:12 exinde quaerebat Pilatus dimittere eum 19:12 And from thenceforth Pilate sought to release him:

but the Jews cried out, saying, If thou let this man go, thou art not Caesar’s friend: whosoever maketh himself a king

Iudaei autem clamabant dicentes si hunc dimittis non es ami- speaketh against Caesar. cus Caesaris omnis qui se regem facit contradicit Caesari 19:13 Pilatus ergo cum audisset hos sermones adduxit foras 19:13 When Pilate therefore heard that saying, he brought

Jesus forth, and sat down in the judgment seat in a place that is called the Pavement, but in the Hebrew, Gabbatha.

Iesum et sedit pro tribunali in locum qui dicitur Lithostrotus hebraice autem Gabbatha 19:14 erat autem parasceue pas- 19:14 And it was the preparation of the passover, and

about the sixth hour: and he saith unto the Jews, Behold your King!

chae hora quasi sexta et dicit Iudaeis ecce rex uester 19:15

19:15 But they cried out, Away with him, away with him, crucify him. Pilate saith unto them, Shall I crucify your King? The chief priests answered, We have no king but Caesar.

illi autem clamabant tolle tolle crucifige eum dixit eis Pilatus regem uestrum crucifigam responderunt pontifices non habemus regem nisi Caesarem 19:16 tunc ergo tradidit eis 19:16 Then delivered he him therefore unto them to be

crucified. And they took Jesus, and led him away.

illum ut crucifigeretur susceperunt autem Iesum et eduxerunt 19:17 et baiulans sibi crucem exiuit in eum qui dicitur 19:17 And he bearing his cross went forth into a place

called the place of a skull, which is called in the Hebrew Golgotha:

Caluariae locum hebraice Golgotha 19:18 ubi eum cruci-

19:18 Where they crucified him, and two other with him, on either side one, and Jesus in the midst.

fixerunt et cum eo alios duos hinc et hinc medium autem Iesum 19:19 scripsit autem et titulum Pilatus et posuit su- 19:19 And Pilate wrote a title, and put it on the cross. And

the writing was JESUS OF NAZARETH THE KING OF THE JEWS.

per crucem erat autem scriptum Iesus Nazarenus rex Iudaeorum 19:20 hunc ergo titulum multi legerunt Iudaeorum quia 19:20 This title then read many of the Jews: for the place

where Jesus was crucified was nigh to the city: and it was written in Hebrew, and Greek, and Latin.

prope ciuitatem erat locus ubi crucifixus est Iesus et erat

Biblia Sacra Vulgata

19:21 Then said the chief priests of the Jews to Pilate, scriptum hebraice graece et latine 19:21 dicebant ergo Pi-

Write not, The King of the Jews; but that he said, I am King of the Jews.

lato pontifices Iudaeorum noli scribere rex Iudaeorum sed

19:22 Pilate answered, What I have written I have written. quia ipse dixit rex sum Iudaeorum 19:22 respondit Pilatus

19:23 Then the soldiers, when they had crucified Jesus, quod scripsi scripsi 19:23 milites ergo cum crucifixissent

took his garments, and made four parts, to every soldier a part; and also his coat: now the coat was without seam,

woven from the top throughout. eum acceperunt uestimenta eius et fecerunt quattuor par-

tes unicuique militi partem et tunicam erat autem tunica in-

19:24 They said therefore among themselves, Let us not consutilis desuper contexta per totum 19:24 dixerunt ergo

rend it, but cast lots for it, whose it shall be: that the scripture might be fulfilled, which saith, They parted my raiment among them, and for my vesture they did cast lots. These things therefore the soldiers did.

ad inuicem non scindamus eam sed sortiamur de illa cuius sit ut scriptura impleatur dicens partiti sunt uestimenta mea sibi et in uestem meam miserunt sortem et milites quidem

19:25 Now there stood by the cross of Jesus his mother, haec fecerunt 19:25 stabant autem iuxta crucem Iesu mater

and his mother’s sister, Mary the wife of Cleophas, and Mary Magdalene.

eius et soror matris eius Maria Cleopae et Maria Magda-

19:26 When Jesus therefore saw his mother, and the disci- lene 19:26 cum uidisset ergo Iesus matrem et discipulum ple standing by, whom he loved, he saith unto his mother, Woman, behold thy son!

stantem quem diligebat dicit matri suae mulier ecce filius

19:27 Then saith he to the disciple, Behold thy mother! tuus 19:27 deinde dicit discipulo ecce mater tua et ex illa

And from that hour that disciple took her unto his own home.

hora accepit eam discipulus in sua 19:28 postea sciens Ie-

19:28 After this, Jesus knowing that all things were now accomplished, that the scripture might be fulfilled, saith, I thirst.

sus quia iam omnia consummata sunt ut consummaretur sc-

19:29 Now there was set a vessel full of vinegar: and they riptura dicit sitio 19:29 uas ergo positum erat aceto plenum

filled a spunge with vinegar, and put it upon hyssop, and put it to his mouth.

illi autem spongiam plenam aceto hysopo circumponentes

19:30 When Jesus therefore had received the vinegar, he obtulerunt ori eius 19:30 cum ergo accepisset Iesus acetum

said, It is finished: and he bowed his head, and gave up the ghost.

dixit consummatum est et inclinato capite tradidit spiritum

19:31 The Jews therefore, because it was the preparation, 19:31 Iudaei ergo quoniam parasceue erat ut non remanerent

that the bodies should not remain upon the cross on the sabbath day, (for that sabbath day was an high day,) besought Pilate that their legs might be broken, and that they might be taken away.

in cruce corpora sabbato erat enim magnus dies ille sabbati rogauerunt Pilatum ut frangerentur eorum crura et tolleren-

19:32 Then came the soldiers, and brake the legs of the tur 19:32 uenerunt ergo milites et primi quidem fregerunt

first, and of the other which was crucified with him.

19:33 But when they came to Jesus, and saw that he was crura et alterius qui crucifixus est cum eo 19:33 ad Iesum

dead already, they brake not his legs:

autem cum uenissent ut uiderunt eum iam mortuum non fre-

19:34 But one of the soldiers with a spear pierced his side, gerunt eius crura 19:34 sed unus militum lancea latus eius and forthwith came there out blood and water.

19:35 And he that saw it bare record, and his record is true: aperuit et continuo exiuit sanguis et aqua 19:35 et qui uidit

and he knoweth that he saith true, that ye might believe.

testimonium perhibuit et uerum est eius testimonium et ille

19:36 For these things were done, that the scripture should scit quia uera dicit ut et uos credatis 19:36 facta sunt enim be fulfilled, A bone of him shall not be broken.

19:37 And again another scripture saith, They shall look haec ut scriptura impleatur os non comminuetis ex eo 19:37

on him whom they pierced.

et iterum alia scriptura dicit uidebunt in quem transfixerunt

Euangelium Secundum Ioannem

19:38 post haec autem rogauit Pilatum Ioseph ab Arimathia 19:38 And after this Joseph of Arimathaea, being a dis-

ciple of Jesus, but secretly for fear of the Jews, besought Pilate that he might take away the body of Jesus: and Pilate gave him leave. He came therefore, and took the body of Jesus.

eo quod esset discipulus Iesu occultus autem propter metum Iudaeorum ut tolleret corpus Iesu et permisit Pilatus uenit ergo et tulit corpus Iesu 19:39 uenit autem et Nicodemus 19:39 And there came also Nicodemus, which at the first

came to Jesus by night, and brought a mixture of myrrh and aloes, about an hundred pound weight.

qui uenerat ad Iesum nocte primum ferens mixturam murrae et aloes quasi libras centum 19:40 acceperunt ergo corpus 19:40 Then took they the body of Jesus, and wound it in

linen clothes with the spices, as the manner of the Jews is to bury.

Iesu et ligauerunt eum linteis cum aromatibus sicut mos Iudaeis est sepelire 19:41 erat autem in loco ubi crucifixus 19:41 Now in the place where he was crucified there was

a garden; and in the garden a new sepulchre, wherein was never man yet laid.

est hortus et in horto monumentum nouum in quo nondum quisquam positus erat 19:42 ibi ergo propter parasceuen Iu- 19:42 There laid they Jesus therefore because of the Jews’

preparation day; for the sepulchre was nigh at hand.

daeorum quia iuxta erat monumentum posuerunt Iesum

20:1 una autem sabbati Maria Magdalene uenit mane cum 20:1 The first day of the week cometh Mary Magdalene

early, when it was yet dark, unto the sepulchre, and seeth the stone taken away from the sepulchre.

adhuc tenebrae essent ad monumentum et uidet lapidem sublatum a monumento 20:2 cucurrit ergo et uenit ad Simo- 20:2 Then she runneth, and cometh to Simon Peter, and to

the other disciple, whom Jesus loved, and saith unto them, They have taken away the LORD out of the sepulchre, and

nem Petrum et ad alium discipulum quem amabat Iesus et we know not where they have laid him. dicit eis tulerunt Dominum de monumento et nescimus ubi posuerunt eum 20:3 exiit ergo Petrus et ille alius discipulus 20:3 Peter therefore went forth, and that other disciple,

and came to the sepulchre.

et uenerunt ad monumentum 20:4 currebant autem duo si- 20:4 So they ran both together: and the other disciple did

outrun Peter, and came first to the sepulchre.

mul et ille alius discipulus praecucurrit citius Petro et uenit

primus ad monumentum 20:5 et cum se inclinasset uidet 20:5 And he stooping down, and looking in, saw the linen

clothes lying; yet went he not in.

posita linteamina non tamen introiuit 20:6 uenit ergo Si- 20:6 Then cometh Simon Peter following him, and went

into the sepulchre, and seeth the linen clothes lie,

mon Petrus sequens eum et introiuit in monumentum et uidet linteamina posita 20:7 et sudarium quod fuerat super caput 20:7 And the napkin, that was about his head, not lying

with the linen clothes, but wrapped together in a place by itself.

eius non cum linteaminibus positum sed separatim inuolutum in unum locum 20:8 tunc ergo introiuit et ille discipu- 20:8 Then went in also that other disciple, which came

first to the sepulchre, and he saw, and believed.

lus qui uenerat primus ad monumentum et uidit et credidit 20:9 nondum enim sciebant scripturam quia oportet eum a 20:9 For as yet they knew not the scripture, that he must

rise again from the dead.

mortuis resurgere 20:10 abierunt ergo iterum ad semet ip- 20:10 Then the disciples went away again unto their own

home.

sos discipuli 20:11 Maria autem stabat ad monumentum fo- 20:11 But Mary stood without at the sepulchre weeping:

and as she wept, she stooped down, and looked into the sepulchre,

ris plorans dum ergo fleret inclinauit se et prospexit in monumentum 20:12 et uidit duos angelos in albis sedentes unum 20:12 And seeth two angels in white sitting, the one at the

head, and the other at the feet, where the body of Jesus had lain.

ad caput et unum ad pedes ubi positum fuerat corpus Iesu 20:13 dicunt ei illi mulier quid ploras dicit eis quia tulerunt 20:13 And they say unto her, Woman, why weepest thou?

She saith unto them, Because they have taken away my LORD, and I know not where they have laid him.

Biblia Sacra Vulgata

20:14 And when she had thus said, she turned herself Dominum meum et nescio ubi posuerunt eum 20:14 haec back, and saw Jesus standing, and knew not that it was Jesus.

cum dixisset conuersa est retrorsum et uidet Iesum stantem

20:15 Jesus saith unto her, Woman, why weepest thou? et non sciebat quia Iesus est 20:15 dicit ei Iesus mulier quid

whom seekest thou? She, supposing him to be the gardener, saith unto him, Sir, if thou have borne him hence,

tell me where thou hast laid him, and I will take him away. ploras quem quaeris illa existimans quia hortulanus esset di-

cit ei domine si tu sustulisti eum dicito mihi ubi posuisti eum

20:16 Jesus saith unto her, Mary. She turned herself, and et ego eum tollam 20:16 dicit ei Iesus Maria conuersa illa saith unto him, Rabboni; which is to say, Master.

20:17 Jesus saith unto her, Touch me not; for I am not yet dicit ei rabboni quod dicitur magister 20:17 dicit ei Iesus

ascended to my Father: but go to my brethren, and say unto them, I ascend unto my Father, and your Father; and

to my God, and your God. noli me tangere nondum enim ascendi ad Patrem meum uade

autem ad fratres meos et dic eis ascendo ad Patrem meum

20:18 Mary Magdalene came and told the disciples that et Patrem uestrum et Deum meum et Deum uestrum 20:18 she had seen the LORD, and that he had spoken these things unto her.

uenit Maria Magdalene adnuntians discipulis quia uidi Do-

20:19 Then the same day at evening, being the first day of minum et haec dixit mihi 20:19 cum esset ergo sero die illo

the week, when the doors were shut where the disciples were assembled for fear of the Jews, came Jesus and stood

in the midst, and saith unto them, Peace be unto you. una sabbatorum et fores essent clausae ubi erant discipuli

propter metum Iudaeorum uenit Iesus et stetit in medio et di-

20:20 And when he had so said, he shewed unto them his cit eis pax uobis 20:20 et hoc cum dixisset ostendit eis ma-

hands and his side. Then were the disciples glad, when they saw the LORD.

nus et latus gauisi sunt ergo discipuli uiso Domino 20:21

20:21 Then said Jesus to them again, Peace be unto you: as my Father hath sent me, even so send I you.

dixit ergo eis iterum pax uobis sicut misit me Pater et ego

20:22 And when he had said this, he breathed on them, mitto uos 20:22 hoc cum dixisset insuflauit et dicit eis ac-

and saith unto them, Receive ye the Holy Ghost:

20:23 Whose soever sins ye remit, they are remitted unto cipite Spiritum Sanctum 20:23 quorum remiseritis peccata

them; and whose soever sins ye retain, they are retained.

20:24 But Thomas, one of the twelve, called Didymus, remittuntur eis quorum retinueritis detenta sunt 20:24 Tho-

was not with them when Jesus came.

mas autem unus ex duodecim qui dicitur Didymus non erat

20:25 The other disciples therefore said unto him, We have cum eis quando uenit Iesus 20:25 dixerunt ergo ei alii disseen the LORD. But he said unto them, Except I shall see in his hands the print of the nails, and put my finger into the print of the nails, and thrust my hand into his side, I will not believe.

cipuli uidimus Dominum ille autem dixit eis nisi uidero in manibus eius figuram clauorum et mittam digitum meum in locum clauorum et mittam manum meam in latus eius non

20:26 And after eight days again his disciples were within, credam 20:26 et post dies octo iterum erant discipuli eius and Thomas with them: then came Jesus, the doors being shut, and stood in the midst, and said, Peace be unto you.

intus et Thomas cum eis uenit Iesus ianuis clausis et stetit

20:27 Then saith he to Thomas, Reach hither thy finger, in medio et dixit pax uobis 20:27 deinde dicit Thomae in-

and behold my hands; and reach hither thy hand, and thrust it into my side: and be not faithless, but believing.

fer digitum tuum huc et uide manus meas et adfer manum tuam et mitte in latus meum et noli esse incredulus sed fi-

20:28 And Thomas answered and said unto him, My delis 20:28 respondit Thomas et dixit ei Dominus meus et

LORD and my God.

20:29 Jesus saith unto him, Thomas, because thou hast Deus meus 20:29 dicit ei Iesus quia uidisti me credidisti

seen me, thou hast believed: blessed are they that have not seen, and yet have believed.

beati qui non uiderunt et crediderunt 20:30 multa quidem

20:30 And many other signs truly did Jesus in the presence of his disciples, which are not written in this book:

Euangelium Secundum Ioannem

et alia signa fecit Iesus in conspectu discipulorum suorum quae non sunt scripta in libro hoc 20:31 haec autem scripta 20:31 But these are written, that ye might believe that Je-

sus is the Christ, the Son of God; and that believing ye might have life through his name.

sunt ut credatis quia Iesus est Christus Filius Dei et ut credentes uitam habeatis in nomine eius

21:1 postea manifestauit se iterum Iesus ad mare Tiberia- 21:1 After these things Jesus shewed himself again to the

disciples at the sea of Tiberias; and on this wise shewed he himself.

dis manifestauit autem sic 21:2 erant simul Simon Petrus et

21:2 There were together Simon Peter, and Thomas called Didymus, and Nathanael of Cana in Galilee, and the sons of Zebedee, and two other of his disciples.

Thomas qui dicitur Didymus et Nathanahel qui erat a Cana Galilaeae et filii Zebedaei et alii ex discipulis eius duo 21:3 21:3 Simon Peter saith unto them, I go a fishing. They

say unto him, We also go with thee. They went forth, and entered into a ship immediately; and that night they caught

dicit eis Simon Petrus uado piscari dicunt ei uenimus et nos nothing. tecum et exierunt et ascenderunt in nauem et illa nocte nihil prendiderunt 21:4 mane autem iam facto stetit Iesus in 21:4 But when the morning was now come, Jesus stood on

the shore: but the disciples knew not that it was Jesus.

litore non tamen cognouerunt discipuli quia Iesus est 21:5 21:5 Then Jesus saith unto them, Children, have ye any

meat? They answered him, No.

dicit ergo eis Iesus pueri numquid pulmentarium habetis responderunt ei non 21:6 dixit eis mittite in dexteram nauigii 21:6 And he said unto them, Cast the net on the right side

of the ship, and ye shall find. They cast therefore, and now they were not able to draw it for the multitude of fishes.

rete et inuenietis miserunt ergo et iam non ualebant illud trahere a multitudine piscium 21:7 dicit ergo discipulus ille 21:7 Therefore that disciple whom Jesus loved saith unto

Peter, It is the Lord. Now when Simon Peter heard that it was the Lord, he girt his fisher’s coat unto him, (for he

quem diligebat Iesus Petro Dominus est Simon Petrus cum was naked,) and did cast himself into the sea. audisset quia Dominus est tunicam succinxit se erat enim

nudus et misit se in mare 21:8 alii autem discipuli nauigio 21:8 And the other disciples came in a little ship; (for they were not far from land, but as it were two hundred cubits,) dragging the net with fishes.

uenerunt non enim longe erant a terra sed quasi a cubitis ducentis trahentes rete piscium 21:9 ut ergo descenderunt 21:9 As soon then as they were come to land, they saw a

fire of coals there, and fish laid thereon, and bread.

in terram uiderunt prunas positas et piscem superpositum et panem 21:10 dicit eis Iesus adferte de piscibus quos pren- 21:10 Jesus saith unto them, Bring of the fish which ye

have now caught.

didistis nunc 21:11 ascendit Simon Petrus et traxit rete in 21:11 Simon Peter went up, and drew the net to land full

of great fishes, an hundred and fifty and three: and for all there were so many, yet was not the net broken.

terram plenum magnis piscibus centum quinquaginta tribus et cum tanti essent non est scissum rete 21:12 dicit eis Iesus 21:12 Jesus saith unto them, Come and dine. And none of

the disciples durst ask him, Who art thou? knowing that it was the Lord.

uenite prandete et nemo audebat discentium interrogare eum tu quis es scientes quia Dominus esset 21:13 et uenit Iesus 21:13 Jesus then cometh, and taketh bread, and giveth

them, and fish likewise.

et accepit panem et dat eis et piscem similiter 21:14 hoc 21:14 This is now the third time that Jesus shewed himself

to his disciples, after that he was risen from the dead.

iam tertio manifestatus est Iesus discipulis cum surrexisset a mortuis 21:15 cum ergo prandissent dicit Simoni Petro Ie- 21:15 So when they had dined, Jesus saith to Simon Peter,

Simon, son of Jonas, lovest thou me more than these? He saith unto him, Yea, Lord; thou knowest that I love thee.

sus Simon Iohannis diligis me plus his dicit ei etiam Domine He saith unto him, Feed my lambs. tu scis quia amo te dicit ei pasce agnos meos 21:16 dicit ei 21:16 He saith to him again the second time, Simon, son

of Jonas, lovest thou me? He saith unto him, Yea, Lord; thou knowest that I love thee. He saith unto him, Feed my sheep.

Biblia Sacra Vulgata

iterum Simon Iohannis diligis me ait illi etiam Domine tu

21:17 He saith unto him the third time, Simon, son of scis quia amo te dicit ei pasce agnos meos 21:17 dicit ei

Jonas, lovest thou me? Peter was grieved because he said unto him the third time, Lovest thou me? And he said unto him, Lord, thou knowest all things; thou knowest that I love thee. Jesus saith unto him, Feed my sheep.

tertio Simon Iohannis amas me contristatus est Petrus quia dixit ei tertio amas me et dicit ei Domine tu omnia scis tu scis

21:18 Verily, verily, I say unto thee, When thou wast quia amo te dicit ei pasce oues meas 21:18 amen amen dico

young, thou girdest thyself, and walkedst whither thou wouldest: but when thou shalt be old, thou shalt stretch forth thy hands, and another shall gird thee, and carry thee whither thou wouldest not.

tibi cum esses iunior cingebas te et ambulabas ubi uolebas cum autem senueris extendes manus tuas et alius te cinget

21:19 This spake he, signifying by what death he should et ducet quo non uis 21:19 hoc autem dixit significans qua

glorify God. And when he had spoken this, he saith unto him, Follow me.

morte clarificaturus esset Deum et hoc cum dixisset dicit ei

21:20 Then Peter, turning about, seeth the disciple whom sequere me 21:20 conuersus Petrus uidit illum discipulum

Jesus loved following; which also leaned on his breast at supper, and said, Lord, which is he that betrayeth thee?

quem diligebat Iesus sequentem qui et recubuit in cena su-

21:21 Peter seeing him saith to Jesus, Lord, and what shall per pectus eius et dixit Domine quis est qui tradit te 21:21 this man do?

hunc ergo cum uidisset Petrus dicit Iesu Domine hic autem

21:22 Jesus saith unto him, If I will that he tarry till I come, quid 21:22 dicit ei Iesus si sic eum uolo manere donec uewhat is that to thee? follow thou me.

21:23 Then went this saying abroad among the brethren, niam quid ad te tu me sequere 21:23 exiuit ergo sermo iste

that that disciple should not die: yet Jesus said not unto him, He shall not die; but, If I will that he tarry till I come,

what is that to thee? in fratres quia discipulus ille non moritur et non dixit ei Iesus

non moritur sed si sic eum uolo manere donec uenio quid ad

21:24 This is the disciple which testifieth of these things, te 21:24 hic est discipulus qui testimonium perhibet de his and wrote these things: and we know that his testimony is true.

et scripsit haec et scimus quia uerum est testimonium eius

21:25 And there are also many other things which Jesus 21:25 sunt autem et alia multa quae fecit Iesus quae si scri-

did, the which, if they should be written every one, I suppose that even the world itself could not contain the books

that should be written. Amen. bantur per singula nec ipsum arbitror mundum capere eos qui scribendi sunt libros amen

Actus Apostolorum

1:1 primum quidem sermonem feci de omnibus o Theo- 1:1 The former treatise have I made, O Theophilus, of all

that Jesus began both to do and teach,

phile quae coepit Iesus facere et docere 1:2 usque in diem 1:2 Until the day in which he was taken up, after that he

through the Holy Ghost had given commandments unto the apostles whom he had chosen:

qua praecipiens apostolis per Spiritum Sanctum quos elegit adsumptus est 1:3 quibus et praebuit se ipsum uiuum post 1:3 To whom also he shewed himself alive after his pas-

sion by many infallible proofs, being seen of them forty days, and speaking of the things pertaining to the kingdom

passionem suam in multis argumentis per dies quadraginta of God: apparens eis et loquens de regno Dei 1:4 et conuescens 1:4 And, being assembled together with them, com-

manded them that they should not depart from Jerusalem, but wait for the promise of the Father, which, saith he, ye

praecepit eis ab Hierosolymis ne discederent sed expecta- have heard of me. rent promissionem Patris quam audistis per os meum 1:5 1:5 For John truly baptized with water; but ye shall be

baptized with the Holy Ghost not many days hence.

quia Iohannes quidem baptizauit aqua uos autem baptizabimini Spiritu Sancto non post multos hos dies 1:6 igitur qui 1:6 When they therefore were come together, they asked

of him, saying, Lord, wilt thou at this time restore again the kingdom to Israel?

conuenerant interrogabant eum dicentes Domine si in tempore hoc restitues regnum Israhel 1:7 dixit autem eis non 1:7 And he said unto them, It is not for you to know the

times or the seasons, which the Father hath put in his own power.

est uestrum nosse tempora uel momenta quae Pater posuit in sua potestate 1:8 sed accipietis uirtutem superuenientis 1:8 But ye shall receive power, after that the Holy Ghost

is come upon you: and ye shall be witnesses unto me both in Jerusalem, and in all Judaea, and in Samaria, and unto

Spiritus Sancti in uos et eritis mihi testes in Hierusalem et the uttermost part of the earth. in omni Iudaea et Samaria et usque ad ultimum terrae 1:9 1:9 And when he had spoken these things, while they be-

held, he was taken up; and a cloud received him out of their sight.

et cum haec dixisset uidentibus illis eleuatus est et nubes

suscepit eum ab oculis eorum 1:10 cumque intuerentur in 1:10 And while they looked stedfastly toward heaven as he

went up, behold, two men stood by them in white apparel;

caelum eunte illo ecce duo uiri adstiterunt iuxta illos in uestibus albis 1:11 qui et dixerunt uiri galilaei quid statis as- 1:11 Which also said, Ye men of Galilee, why stand ye

gazing up into heaven? this same Jesus, which is taken up from you into heaven, shall so come in like manner as ye

picientes in caelum hic Iesus qui adsumptus est a uobis in have seen him go into heaven. caelum sic ueniet quemadmodum uidistis eum euntem in caelum 1:12 tunc reuersi sunt Hierosolymam a monte qui 1:12 Then returned they unto Jerusalem from the mount

called Olivet, which is from Jerusalem a sabbath day’s journey.

uocatur Oliueti qui est iuxta Hierusalem sabbati habens iter 1:13 et cum introissent in cenaculum ascenderunt ubi ma- 1:13 And when they were come in, they went up into

an upper room, where abode both Peter, and James, and John, and Andrew, Philip, and Thomas, Bartholomew, and Matthew, James the son of Alphaeus, and Simon Zelotes, and Judas the brother of James.

nebant Petrus et Iohannes Iacobus et Andreas Philippus et Thomas Bartholomeus et Mattheus Iacobus Alphei et Simon Zelotes et Iudas Iacobi 1:14 hii omnes erant perseuerantes 1:14 These all continued with one accord in prayer and

supplication, with the women, and Mary the mother of Jesus, and with his brethren.

unianimiter in oratione cum mulieribus et Maria matre Iesu et fratribus eius 1:15 et in diebus illis exsurgens Petrus in 1:15 And in those days Peter stood up in the midst of the

disciples, and said, (the number of names together were about an hundred and twenty,)

medio fratrum dixit erat autem turba nominum simul fere centum uiginti 1:16 uiri fratres oportet impleri scripturam 1:16 Men and brethren, this scripture must needs have

been fulfilled, which the Holy Ghost by the mouth of David spake before concerning Judas, which was guide

quam praedixit Spiritus Sanctus per os Dauid de Iuda qui to them that took Jesus.

Biblia Sacra Vulgata

1:17 For he was numbered with us, and had obtained part fuit dux eorum qui conprehenderunt Iesum 1:17 quia con-

of this ministry.

numeratus erat in nobis et sortitus est sortem ministerii huius

1:18 Now this man purchased a field with the reward of in- 1:18 et hic quidem possedit agrum de mercede iniquitatis et

iquity; and falling headlong, he burst asunder in the midst, and all his bowels gushed out.

suspensus crepuit medius et diffusa sunt omnia uiscera eius

1:19 And it was known unto all the dwellers at Jerusalem; 1:19 et notum factum est omnibus habitantibus Hierusalem

insomuch as that field is called in their proper tongue, Aceldama, that is to say, The field of blood.

ita ut appellaretur ager ille lingua eorum Acheldemach hoc

1:20 For it is written in the book of Psalms, Let his habi- est ager Sanguinis 1:20 scriptum est enim in libro Psalmo-

tation be desolate, and let no man dwell therein: and his bishoprick let another take.

rum fiat commoratio eius deserta et non sit qui inhabitet in ea

1:21 Wherefore of these men which have companied with et episcopatum eius accipiat alius 1:21 oportet ergo ex his

us all the time that the Lord Jesus went in and out among us,

uiris qui nobiscum congregati sunt in omni tempore quo in-

1:22 Beginning from the baptism of John, unto that same trauit et exiuit inter nos Dominus Iesus 1:22 incipiens a ba-

day that he was taken up from us, must one be ordained to be a witness with us of his resurrection.

ptismate Iohannis usque in diem qua adsumptus est a nobis

1:23 And they appointed two, Joseph called Barsabas, testem resurrectionis eius nobiscum fieri unum ex istis 1:23

who was surnamed Justus, and Matthias.

et statuerunt duos Ioseph qui uocabatur Barsabban qui cog-

1:24 And they prayed, and said, Thou, Lord, which know- nominatus est Iustus et Matthiam 1:24 et orantes dixerunt

est the hearts of all men, shew whether of these two thou hast chosen,

tu Domine qui corda nosti omnium ostende quem elegeris

1:25 That he may take part of this ministry and apostle- ex his duobus unum 1:25 accipere locum ministerii huius

ship, from which Judas by transgression fell, that he might go to his own place.

et apostolatus de quo praeuaricatus est Iudas ut abiret in lo-

1:26 And they gave forth their lots; and the lot fell upon cum suum 1:26 et dederunt sortes eis et cecidit sors super

Matthias; and he was numbered with the eleven apostles.

Matthiam et adnumeratus est cum undecim apostolis

2:1 And when the day of Pentecost was fully come, they 2:1 et cum conplerentur dies pentecostes erant omnes pa-

were all with one accord in one place.

2:2 And suddenly there came a sound from heaven as of a riter in eodem loco 2:2 et factus est repente de caelo so-

rushing mighty wind, and it filled all the house where they were sitting.

nus tamquam aduenientis spiritus uehementis et repleuit to-

2:3 And there appeared unto them cloven tongues like as tam domum ubi erant sedentes 2:3 et apparuerunt illis dis-

of fire, and it sat upon each of them.

pertitae linguae tamquam ignis seditque supra singulos eo-

2:4 And they were all filled with the Holy Ghost, and be- rum 2:4 et repleti sunt omnes Spiritu Sancto et coeperunt

gan to speak with other tongues, as the Spirit gave them utterance.

loqui aliis linguis prout Spiritus Sanctus dabat eloqui illis

2:5 And there were dwelling at Jerusalem Jews, devout 2:5 erant autem in Hierusalem habitantes Iudaei uiri religiosi

men, out of every nation under heaven. 2:6 Now when this was noised abroad, the multitude came ex omni natione quae sub caelo sunt 2:6 facta autem hac together, and were confounded, because that every man heard them speak in his own language.

uoce conuenit multitudo et mente confusa est quoniam au-

2:7 And they were all amazed and marvelled, saying diebat unusquisque lingua sua illos loquentes 2:7 stupebant

one to another, Behold, are not all these which speak Galilaeans?

autem omnes et mirabantur dicentes nonne omnes ecce isti

2:8 And how hear we every man in our own tongue, qui loquuntur Galilaei sunt 2:8 et quomodo nos audiuimus

wherein we were born?

Actus Apostolorum

unusquisque lingua nostra in qua nati sumus 2:9 Parthi et 2:9 Parthians, and Medes, and Elamites, and the dwellers

in Mesopotamia, and in Judaea, and Cappadocia, in Pontus, and Asia,

Medi et Elamitae et qui habitant Mesopotamiam et Iudaeam

et Cappadociam Pontum et Asiam 2:10 Frygiam et Pam- 2:10 Phrygia, and Pamphylia, in Egypt, and in the parts of Libya about Cyrene, and strangers of Rome, Jews and proselytes,

philiam Aegyptum et partes Lybiae quae est circa Cyrenen et aduenae romani 2:11 Iudaei quoque et proselyti Cretes 2:11 Cretes and Arabians, we do hear them speak in our

tongues the wonderful works of God.

et Arabes audiuimus loquentes eos nostris linguis magnalia Dei 2:12 stupebant autem omnes et mirabantur ad in- 2:12 And they were all amazed, and were in doubt, saying

one to another, What meaneth this?

uicem dicentes quidnam hoc uult esse 2:13 alii autem in- 2:13 Others mocking said, These men are full of new wine. ridentes dicebant quia musto pleni sunt isti 2:14 stans au- 2:14 But Peter, standing up with the eleven, lifted up his

voice, and said unto them, Ye men of Judaea, and all ye that dwell at Jerusalem, be this known unto you, and hear-

tem Petrus cum undecim leuauit uocem suam et locutus est ken to my words: eis uiri iudaei et qui habitatis Hierusalem uniuersi hoc uobis notum sit et auribus percipite uerba mea 2:15 non enim sic- 2:15 For these are not drunken, as ye suppose, seeing it is

but the third hour of the day.

ut uos aestimatis hii ebrii sunt cum sit hora diei tertia 2:16 2:16 But this is that which was spoken by the prophet Joel;

sed hoc est quod dictum est per prophetam Iohel 2:17 et 2:17 And it shall come to pass in the last days, saith God, I will pour out of my Spirit upon all flesh: and your sons and your daughters shall prophesy, and your young men

erit in nouissimis diebus dicit Dominus effundam de Spi- shall see visions, and your old men shall dream dreams: ritu meo super omnem carnem et prophetabunt filii uestri et filiae uestrae et iuuenes uestri uisiones uidebunt et seniores uestri somnia somniabunt 2:18 et quidem super seruos 2:18 And on my servants and on my handmaidens I will

pour out in those days of my Spirit; and they shall prophesy:

meos et super ancillas meas in diebus illis effundam de Spiritu meo et prophetabunt 2:19 et dabo prodiiia in caelo sur- 2:19 And I will shew wonders in heaven above, and signs

in the earth beneath; blood, and fire, and vapour of smoke:

sum et signa in terra deorsum sanguinem et ignem et uaporem fumi 2:20 sol conuertetur in tenebras et luna in san- 2:20 The sun shall be turned into darkness, and the moon

into blood, before the great and notable day of the Lord come:

guinem antequam ueniat dies Domini magnus et manifes-

tus 2:21 et erit omnis quicumque inuocauerit nomen Do- 2:21 And it shall come to pass, that whosoever shall call

on the name of the Lord shall be saved.

mini saluus erit 2:22 uiri israhelitae audite uerba haec Ie- 2:22 Ye men of Israel, hear these words; Jesus of Nazareth,

a man approved of God among you by miracles and wonders and signs, which God did by him in the midst of you,

sum Nazarenum uirum adprobatum a Deo in uobis uirtuti- as ye yourselves also know: bus et prodiiiis et signis quae fecit per illum Deus in medio uestri sicut uos scitis 2:23 hunc definito consilio et prae- 2:23 Him, being delivered by the determinate counsel and

foreknowledge of God, ye have taken, and by wicked hands have crucified and slain:

scientia Dei traditum per manus iniquorum adfigentes interemistis 2:24 quem Deus suscitauit solutis doloribus inferni 2:24 Whom God hath raised up, having loosed the pains

of death: because it was not possible that he should be holden of it.

iuxta quod inpossibile erat teneri illum ab eo 2:25 Dauid

2:25 For David speaketh concerning him, I foresaw the Lord always before my face, for he is on my right hand, that I should not be moved:

enim dicit in eum prouidebam Dominum coram me semper quoniam a dextris meis est ne commouear 2:26 propter hoc 2:26 Therefore did my heart rejoice, and my tongue was

glad; moreover also my flesh shall rest in hope:

Biblia Sacra Vulgata

laetatum est cor meum et exultauit lingua mea insuper et

2:27 Because thou wilt not leave my soul in hell, neither caro mea requiescet in spe 2:27 quoniam non derelinques wilt thou suffer thine Holy One to see corruption.

animam meam in inferno neque dabis Sanctum tuum uidere

2:28 Thou hast made known to me the ways of life; thou corruptionem 2:28 notas fecisti mihi uias uitae replebis me

shalt make me full of joy with thy countenance.

2:29 Men and brethren, let me freely speak unto you of the iucunditate cum facie tua 2:29 uiri fratres liceat audenter

patriarch David, that he is both dead and buried, and his sepulchre is with us unto this day.

dicere ad uos de patriarcha Dauid quoniam et defunctus est et sepultus est et sepulchrum eius est apud nos usque in ho-

2:30 Therefore being a prophet, and knowing that God had diernum diem 2:30 propheta igitur cum esset et sciret quia

sworn with an oath to him, that of the fruit of his loins, according to the flesh, he would raise up Christ to sit on

his throne; iureiurando iurasset illi Deus de fructu lumbi eius sedere su-2:31 He seeing this before spake of the resurrection of per sedem eius 2:31 prouidens locutus est de resurrectione

Christ, that his soul was not left in hell, neither his flesh did see corruption.

Christi quia neque derelictus est in inferno neque caro eius

2:32 This Jesus hath God raised up, whereof we all are uidit corruptionem 2:32 hunc Iesum resuscitauit Deus cui

witnesses.

2:33 Therefore being by the right hand of God exalted, omnes nos testes sumus 2:33 dextera igitur Dei exaltatus

and having received of the Father the promise of the Holy Ghost, he hath shed forth this, which ye now see and hear.

et promissione Spiritus Sancti accepta a Patre effudit hunc

2:34 For David is not ascended into the heavens: but he quem uos uidetis et audistis 2:34 non enim Dauid ascendit

saith himself, The Lord said unto my Lord, Sit thou on my right hand,

in caelos dicit autem ipse dixit Dominus Domino meo sede a

2:35 Until I make thy foes thy footstool. dextris meis 2:35 donec ponam inimicos tuos scabillum pe-

2:36 Therefore let all the house of Israel know assuredly, dum tuorum 2:36 certissime ergo sciat omnis domus Israhel that God hath made the same Jesus, whom ye have crucified, both Lord and Christ.

quia et Dominum eum et Christum Deus fecit hunc Iesum

2:37 Now when they heard this, they were pricked in their quem uos crucifixistis 2:37 his auditis conpuncti sunt corde

heart, and said unto Peter and to the rest of the apostles, Men and brethren, what shall we do?

et dixerunt ad Petrum et ad reliquos apostolos quid facie-

2:38 Then Peter said unto them, Repent, and be baptized mus uiri fratres 2:38 Petrus uero ad illos paenitentiam in-

every one of you in the name of Jesus Christ for the remission of sins, and ye shall receive the gift of the Holy

Ghost. quit agite et baptizetur unusquisque uestrum in nomine Iesu

Christi in remissionem peccatorum uestrorum et accipietis

2:39 For the promise is unto you, and to your children, and donum Sancti Spiritus 2:39 uobis enim est repromissio et to all that are afar off, even as many as the LORD our God shall call.

filiis uestris et omnibus qui longe sunt quoscumque aduo-

2:40 And with many other words did he testify and exhort, cauerit Dominus Deus noster 2:40 aliis etiam uerbis plu-

saying, Save yourselves from this untoward generation.

ribus testificatus est et exhortabatur eos dicens saluamini a

2:41 Then they that gladly received his word were bap- generatione ista praua 2:41 qui ergo receperunt sermonem

tized: and the same day there were added unto them about three thousand souls.

eius baptizati sunt et adpositae sunt in illa die animae cir-

2:42 And they continued stedfastly in the apostles’ doc- citer tria milia 2:42 erant autem perseuerantes in doctrina

trine and fellowship, and in breaking of bread, and in prayers.

apostolorum et communicatione fractionis panis et oratio-

2:43 And fear came upon every soul: and many wonders nibus 2:43 fiebat autem omni animae timor multa quoque

and signs were done by the apostles.

Actus Apostolorum

prodiiia et signa per apostolos fiebant in Hierusalem et me-

tus erat magnus in uniuersis 2:44 omnes etiam qui crede- 2:44 And all that believed were together, and had all things

common;

bant erant pariter et habebant omnia communia 2:45 pos- 2:45 And sold their possessions and goods, and parted

them to all men, as every man had need.

sessiones et substantias uendebant et diuidebant illa omnibus prout cuique opus erat 2:46 cotidie quoque perdurantes 2:46 And they, continuing daily with one accord in the

temple, and breaking bread from house to house, did eat their meat with gladness and singleness of heart,

unianimiter in templo et frangentes circa domos panem sumebant cibum cum exultatione et simplicitate cordis 2:47 2:47 Praising God, and having favour with all the people.

And the Lord added to the church daily such as should be saved.

conlaudantes Deum et habentes gratiam ad omnem plebem Dominus autem augebat qui salui fierent cotidie in id ipsum

3:1 Petrus autem et Iohannes ascendebant in templum ad 3:1 Now Peter and John went up together into the temple

at the hour of prayer, being the ninth hour.

horam orationis nonam 3:2 et quidam uir qui erat claudus 3:2 And a certain man lame from his mother’s womb was

carried, whom they laid daily at the gate of the temple which is called Beautiful, to ask alms of them that entered

ex utero matris suae baiulabatur quem ponebant cotidie ad into the temple; portam templi quae dicitur Speciosa ut peteret elemosynam ab introeuntibus in templum 3:3 is cum uidisset Petrum et 3:3 Who seeing Peter and John about to go into the temple

asked an alms.

Iohannem incipientes introire in templum rogabat ut elemo-

synam acciperet 3:4 intuens autem in eum Petrus cum Io- 3:4 And Peter, fastening his eyes upon him with John, said,

Look on us.

hanne dixit respice in nos 3:5 at ille intendebat in eos spe- 3:5 And he gave heed unto them, expecting to receive

something of them.

rans se aliquid accepturum ab eis 3:6 Petrus autem dixit 3:6 Then Peter said, Silver and gold have I none; but

such as I have give I thee: In the name of Jesus Christ of Nazareth rise up and walk.

argentum et aurum non est mihi quod autem habeo hoc tibi do in nomine Iesu Christi Nazareni surge et ambula 3:7 et 3:7 And he took him by the right hand, and lifted him

up: and immediately his feet and ankle bones received strength.

adprehensa ei manu dextera adleuauit eum et protinus consolidatae sunt bases eius et plantae 3:8 et exiliens stetit et 3:8 And he leaping up stood, and walked, and entered with

them into the temple, walking, and leaping, and praising God.

ambulabat et intrauit cum illis in templum ambulans et exiliens et laudans Dominum 3:9 et uidit omnis populus eum 3:9 And all the people saw him walking and praising God: ambulantem et laudantem Deum 3:10 cognoscebant autem 3:10 And they knew that it was he which sat for alms at

the Beautiful gate of the temple: and they were filled with wonder and amazement at that which had happened unto

illum quoniam ipse erat qui ad elemosynam sedebat ad Spe- him. ciosam portam templi et impleti sunt stupore et extasi in eo quod contigerat illi 3:11 cum teneret autem Petrum et Io- 3:11 And as the lame man which was healed held Peter

and John, all the people ran together unto them in the porch that is called Solomon’s, greatly wondering.

hannem concurrit omnis populus ad eos ad porticum qui appellatur Salomonis stupentes 3:12 uidens autem Petrus res- 3:12 And when Peter saw it, he answered unto the people,

Ye men of Israel, why marvel ye at this? or why look ye so earnestly on us, as though by our own power or holiness

pondit ad populum uiri israhelitae quid miramini in hoc aut we had made this man to walk? nos quid intuemini quasi nostra uirtute aut pietate fecerimus hunc ambulare 3:13 Deus Abraham et Deus Isaac et Deus 3:13 The God of Abraham, and of Isaac, and of Jacob,

the God of our fathers, hath glorified his Son Jesus; whom ye delivered up, and denied him in the presence of Pilate, when he was determined to let him go.

Biblia Sacra Vulgata

Iacob Deus patrum nostrorum glorificauit Filium suum Iesum quem uos quidem tradidistis et negastis ante faciem Pi-

3:14 But ye denied the Holy One and the Just, and desired lati iudicante illo dimitti 3:14 uos autem sanctum et iustum

a murderer to be granted unto you;

3:15 And killed the Prince of life, whom God hath raised negastis et petistis uirum homicidam donari uobis 3:15 auc-

from the dead; whereof we are witnesses.

torem uero uitae interfecistis quem Deus suscitauit a mortuis

3:16 And his name through faith in his name hath made cuius nos testes sumus 3:16 et in fide nominis eius hunc

this man strong, whom ye see and know: yea, the faith which is by him hath given him this perfect soundness in

the presence of you all. quem uidetis et nostis confirmauit nomen eius et fides quae

per eum est dedit integram sanitatem istam in conspectu om-

3:17 And now, brethren, I wot that through ignorance ye nium uestrum 3:17 et nunc fratres scio quia per ignorantiam

did it, as did also your rulers.

3:18 But those things, which God before had shewed by fecistis sicut et principes uestri 3:18 Deus autem quae prae-

the mouth of all his prophets, that Christ should suffer, he hath so fulfilled.

nuntiauit per os omnium prophetarum pati Christum suum

3:19 Repent ye therefore, and be converted, that your sins impleuit sic 3:19 paenitemini igitur et conuertimini ut de-

may be blotted out, when the times of refreshing shall come from the presence of the Lord.

leantur uestra peccata 3:20 ut cum uenerint tempora refri-

3:20 And he shall send Jesus Christ, which before was preached unto you:

gerii a conspectu Domini et miserit eum qui praedicatus est

3:21 Whom the heaven must receive until the times of uobis Iesum Christum 3:21 quem oportet caelum quidem

restitution of all things, which God hath spoken by the mouth of all his holy prophets since the world began.

suscipere usque in tempora restitutionis omnium quae locutus est Deus per os sanctorum suorum a saeculo prophetarum

3:22 For Moses truly said unto the fathers, A prophet shall 3:22 Moses quidem dixit quia prophetam uobis suscitabit Do-

the Lord your God raise up unto you of your brethren, like unto me; him shall ye hear in all things whatsoever he shall

say unto you. minus Deus uester de fratribus uestris tamquam me ipsum 3:23 And it shall come to pass, that every soul, which will audietis iuxta omnia quaecumque locutus fuerit uobis 3:23

not hear that prophet, shall be destroyed from among the people.

erit autem omnis anima quae non audierit prophetam illum

3:24 Yea, and all the prophets from Samuel and those that exterminabitur de plebe 3:24 et omnes prophetae a Samufollow after, as many as have spoken, have likewise foretold of these days.

hel et deinceps qui locuti sunt et adnuntiauerunt dies istos

3:25 Ye are the children of the prophets, and of the 3:25 uos estis filii prophetarum et testamenti quod disposuit

covenant which God made with our fathers, saying unto Abraham, And in thy seed shall all the kindreds of the

earth be blessed. Deus ad patres uestros dicens ad Abraham et in semine tuo 3:26 Unto you first God, having raised up his Son Jesus, benedicentur omnes familiae terrae 3:26 uobis primum De-

sent him to bless you, in turning away every one of you from his iniquities.

us suscitans Filium suum misit eum benedicentem uobis ut conuertat se unusquisque a nequitia sua

4:1 And as they spake unto the people, the priests, and 4:1 loquentibus autem illis ad populum superuenerunt sa-

the captain of the temple, and the Sadducees, came upon them,

cerdotes et magistratus templi et Sadducaei 4:2 dolentes

4:2 Being grieved that they taught the people, and preached through Jesus the resurrection from the dead.

quod docerent populum et adnuntiarent in Iesu resurrectio-

4:3 And they laid hands on them, and put them in hold nem ex mortuis 4:3 et iniecerunt in eis manus et posuerunt

unto the next day: for it was now eventide. 4:4 Howbeit many of them which heard the word believed; eos in custodiam in crastinum erat enim iam uespera 4:4 and the number of the men was about five thousand.

Actus Apostolorum

multi autem eorum qui audierant uerbum crediderunt et factus est numerus uirorum quinque milia 4:5 factum est au- 4:5 And it came to pass on the morrow, that their rulers,

and elders, and scribes,

tem in crastinum ut congregarentur principes eorum et seniores et scribae in Hierusalem 4:6 et Annas princeps sacerdo- 4:6 And Annas the high priest, and Caiaphas, and John,

and Alexander, and as many as were of the kindred of the high priest, were gathered together at Jerusalem.

tum et Caiphas et Iohannes et Alexander et quotquot erant de genere sacerdotali 4:7 et statuentes eos in medio interroga- 4:7 And when they had set them in the midst, they asked,

By what power, or by what name, have ye done this?

bant in qua uirtute aut in quo nomine fecistis hoc uos 4:8 4:8 Then Peter, filled with the Holy Ghost, said unto them,

Ye rulers of the people, and elders of Israel,

tunc Petrus repletus Spiritu Sancto dixit ad eos principes populi et seniores 4:9 si nos hodie diiudicamur in benefacto 4:9 If we this day be examined of the good deed done to

the impotent man, by what means he is made whole;

hominis infirmi in quo iste saluus factus est 4:10 notum sit 4:10 Be it known unto you all, and to all the people of

Israel, that by the name of Jesus Christ of Nazareth, whom ye crucified, whom God raised from the dead, even by him

omnibus uobis et omni plebi Israhel quia in nomine Iesu Ch- doth this man stand here before you whole. risti Nazareni quem uos crucifixistis quem Deus suscitauit a mortuis in hoc iste adstat coram uobis sanus 4:11 hic est 4:11 This is the stone which was set at nought of you

builders, which is become the head of the corner.

lapis qui reprobatus est a uobis aedificantibus qui factus est in caput anguli 4:12 et non est in alio aliquo salus nec enim 4:12 Neither is there salvation in any other: for there is

none other name under heaven given among men, whereby we must be saved.

nomen aliud est sub caelo datum hominibus in quo oportet nos saluos fieri 4:13 uidentes autem Petri constantiam et Io- 4:13 Now when they saw the boldness of Peter and John,

and perceived that they were unlearned and ignorant men, they marvelled; and they took knowledge of them, that

hannis conperto quod homines essent sine litteris et idiotae they had been with Jesus. admirabantur et cognoscebant eos quoniam cum Iesu fuerant 4:14 hominem quoque uidentes stantem cum eis qui cura- 4:14 And beholding the man which was healed standing

with them, they could say nothing against it.

tus fuerat nihil poterant contradicere 4:15 iusserunt autem 4:15 But when they had commanded them to go aside out

of the council, they conferred among themselves,

eos foras extra concilium secedere et conferebant ad inuicem 4:16 dicentes quid faciemus hominibus istis quoniam quidem 4:16 Saying, What shall we do to these men? for that in-

deed a notable miracle hath been done by them is manifest to all them that dwell in Jerusalem; and we cannot deny it.

notum signum factum est per eos omnibus habitantibus in Hierusalem manifestum et non possumus negare 4:17 sed 4:17 But that it spread no further among the people, let

us straitly threaten them, that they speak henceforth to no man in this name.

ne amplius diuulgetur in populum comminemur eis ne ultra loquantur in nomine hoc ulli hominum 4:18 et uocantes 4:18 And they called them, and commanded them not to

speak at all nor teach in the name of Jesus.

eos denuntiauerunt ne omnino loquerentur neque docerent in nomine Iesu 4:19 Petrus uero et Iohannes respondentes 4:19 But Peter and John answered and said unto them,

Whether it be right in the sight of God to hearken unto you more than unto God, judge ye.

dixerunt ad eos si iustum est in conspectu Dei uos potius audire quam Deum iudicate 4:20 non enim possumus quae 4:20 For we cannot but speak the things which we have

seen and heard.

uidimus et audiuimus non loqui 4:21 at illi comminantes di- 4:21 So when they had further threatened them, they let

them go, finding nothing how they might punish them, because of the people: for all men glorified God for that

miserunt eos non inuenientes quomodo punirent eos propter which was done.

Biblia Sacra Vulgata

populum quia omnes clarificabant Deum in eo quod accide-

4:22 For the man was above forty years old, on whom this rat 4:22 annorum enim erat amplius quadraginta homo in

miracle of healing was shewed.

4:23 And being let go, they went to their own company, quo factum erat signum istud sanitatis 4:23 dimissi autem

and reported all that the chief priests and elders had said unto them.

uenerunt ad suos et adnuntiauerunt eis quanta ad eos prin-

4:24 And when they heard that, they lifted up their voice to cipes sacerdotum et seniores dixissent 4:24 qui cum audis-

God with one accord, and said, Lord, thou art God, which hast made heaven, and earth, and the sea, and all that in

them is: sent unianimiter leuauerunt uocem ad Deum et dixerunt Do-

mine tu qui fecisti caelum et terram et mare et omnia quae

4:25 Who by the mouth of thy servant David hast said, in eis sunt 4:25 qui Spiritu Sancto per os patris nostri Da-

Why did the heathen rage, and the people imagine vain things?

uid pueri tui dixisti quare fremuerunt gentes et populi me-

4:26 The kings of the earth stood up, and the rulers were ditati sunt inania 4:26 adstiterunt reges terrae et principes

gathered together against the Lord, and against his Christ.

conuenerunt in unum aduersus Dominum et aduersus Chri-

4:27 For of a truth against thy holy child Jesus, whom stum eius 4:27 conuenerunt enim uere in ciuitate ista ad-

thou hast anointed, both Herod, and Pontius Pilate, with the Gentiles, and the people of Israel, were gathered to-

gether, uersus sanctum puerum tuum Iesum quem unxisti Herodes 4:28 For to do whatsoever thy hand and thy counsel deter- et Pontius Pilatus cum gentibus et populis Israhel 4:28 fa-

mined before to be done. 4:29 And now, Lord, behold their threatenings: and grant cere quae manus tua et consilium decreuerunt fieri 4:29 et unto thy servants, that with all boldness they may speak thy word,

nunc Domine respice in minas eorum et da seruis tuis cum

4:30 By stretching forth thine hand to heal; and that signs omni fiducia loqui uerbum tuum 4:30 in eo cum manum and wonders may be done by the name of thy holy child Jesus.

tuam extendas sanitates et signa et prodiiia fieri per nomen

4:31 And when they had prayed, the place was shaken sancti Filii tui Iesu 4:31 et cum orassent motus est locus

where they were assembled together; and they were all filled with the Holy Ghost, and they spake the word of

God with boldness. in quo erant congregati et repleti sunt omnes Spiritu Sancto 4:32 And the multitude of them that believed were of one et loquebantur uerbum Dei cum fiducia 4:32 multitudinis

heart and of one soul: neither said any of them that ought of the things which he possessed was his own; but they

had all things common. autem credentium erat cor et anima una nec quisquam eo-

rum quae possidebant aliquid suum esse dicebat sed erant

4:33 And with great power gave the apostles witness of the illis omnia communia 4:33 et uirtute magna reddebant apo-

resurrection of the Lord Jesus: and great grace was upon them all.

stoli testimonium resurrectionis Iesu Christi Domini et gra-

4:34 Neither was there any among them that lacked: for tia magna erat in omnibus illis 4:34 neque enim quisquam

as many as were possessors of lands or houses sold them, and brought the prices of the things that were sold,

egens erat inter illos quotquot enim possessores agrorum aut domorum erant uendentes adferebant pretia eorum quae uen-

4:35 And laid them down at the apostles’ feet: and dis- debant 4:35 et ponebant ante pedes apostolorum diuideban-

tribution was made unto every man according as he had need.

tur autem singulis prout cuique opus erat 4:36 Ioseph autem

4:36 And Joses, who by the apostles was surnamed Barnabas, (which is, being interpreted, The son of consolation,) a Levite, and of the country of Cyprus,

qui cognominatus est Barnabas ab apostolis quod est inter-

4:37 Having land, sold it, and brought the money, and laid pretatum Filius consolationis Leuites Cyprius genere 4:37

it at the apostles’ feet.

cum haberet agrum uendidit illum et adtulit pretium et posuit

Actus Apostolorum

ante pedes apostolorum

5:1 uir autem quidam nomine Ananias cum Saffira uxore 5:1 But a certain man named Ananias, with Sapphira his

wife, sold a possession,

sua uendidit agrum 5:2 et fraudauit de pretio agri conscia 5:2 And kept back part of the price, his wife also being

privy to it, and brought a certain part, and laid it at the apostles’ feet.

uxore sua et adferens partem quandam ad pedes apostolorum posuit 5:3 dixit autem Petrus Anania cur temptauit Satanas 5:3 But Peter said, Ananias, why hath Satan filled thine

heart to lie to the Holy Ghost, and to keep back part of the price of the land?

cor tuum mentiri te Spiritui Sancto et fraudare de pretio agri 5:4 nonne manens tibi manebat et uenundatum in tua erat po- 5:4 Whiles it remained, was it not thine own? and after it

was sold, was it not in thine own power? why hast thou conceived this thing in thine heart? thou hast not lied unto

testate quare posuisti in corde tuo hanc rem non es mentitus men, but unto God. hominibus sed Deo 5:5 audiens autem Ananias haec uerba 5:5 And Ananias hearing these words fell down, and gave

up the ghost: and great fear came on all them that heard these things.

cecidit et exspirauit et factus est timor magnus in omnes qui audierant 5:6 surgentes autem iuuenes amouerunt eum et 5:6 And the young men arose, wound him up, and carried

him out, and buried him.

efferentes sepelierunt 5:7 factum est autem quasi horarum 5:7 And it was about the space of three hours after, when

his wife, not knowing what was done, came in.

trium spatium et uxor ipsius nesciens quod factum fuerat introiit 5:8 respondit autem ei Petrus dic mihi si tanti agrum 5:8 And Peter answered unto her, Tell me whether ye sold

the land for so much? And she said, Yea, for so much.

uendidistis at illa dixit etiam tanti 5:9 Petrus autem ad eam 5:9 Then Peter said unto her, How is it that ye have agreed

together to tempt the Spirit of the Lord? behold, the feet of them which have buried thy husband are at the door, and

quid utique conuenit uobis temptare Spiritum Domini ecce shall carry thee out. pedes eorum qui sepelierunt uirum tuum ad ostium et efferent te 5:10 confestim cecidit ante pedes eius et exspirauit 5:10 Then fell she down straightway at his feet, and

yielded up the ghost: and the young men came in, and found her dead, and, carrying her forth, buried her by her

intrantes autem iuuenes inuenerunt illam mortuam et extule- husband. runt et sepelierunt ad uirum suum 5:11 et factus est timor 5:11 And great fear came upon all the church, and upon as

many as heard these things.

magnus in uniuersa ecclesia et in omnes qui audierunt haec 5:12 per manus autem apostolorum fiebant signa et prodiiia 5:12 And by the hands of the apostles were many signs

and wonders wrought among the people; (and they were all with one accord in Solomon’s porch.

multa in plebe et erant unianimiter omnes in porticu Salomonis 5:13 ceterorum autem nemo audebat coniungere se 5:13 And of the rest durst no man join himself to them:

but the people magnified them.

illis sed magnificabat eos populus 5:14 magis autem auge- 5:14 And believers were the more added to the Lord, mul-

titudes both of men and women.)

batur credentium in Domino multitudo uirorum ac mulierum 5:15 ita ut in plateas eicerent infirmos et ponerent in lectulis 5:15 Insomuch that they brought forth the sick into the

streets, and laid them on beds and couches, that at the least the shadow of Peter passing by might overshadow some of

et grabattis ut ueniente Petro saltim umbra illius obumbra- them. ret quemquam eorum 5:16 concurrebat autem et multitudo 5:16 There came also a multitude out of the cities round

about unto Jerusalem, bringing sick folks, and them which were vexed with unclean spirits: and they were healed ev-

uicinarum ciuitatum Hierusalem adferentes aegros et uexa- ery one. tos ab spiritibus inmundis qui curabantur omnes 5:17 ex- 5:17 Then the high priest rose up, and all they that were

with him, (which is the sect of the Sadducees,) and were filled with indignation,

surgens autem princeps sacerdotum et omnes qui cum illo erant quae est heresis Sadducaeorum repleti sunt zelo 5:18 5:18 And laid their hands on the apostles, and put them in

the common prison.

Biblia Sacra Vulgata

et iniecerunt manus in apostolos et posuerunt illos in cus-

5:19 But the angel of the Lord by night opened the prison todia publica 5:19 angelus autem Domini per noctem ape-

doors, and brought them forth, and said,

5:20 Go, stand and speak in the temple to the people all riens ianuas carceris et educens eos dixit 5:20 ite et stantes

the words of this life.

5:21 And when they heard that, they entered into the tem- loquimini in templo plebi omnia uerba uitae huius 5:21 qui

ple early in the morning, and taught. But the high priest came, and they that were with him, and called the council together, and all the senate of the children of Israel, and sent to the prison to have them brought.

cum audissent intrauerunt diluculo in templum et docebant adueniens autem princeps sacerdotum et qui cum eo erant conuocauerunt concilium et omnes seniores filiorum Israhel

5:22 But when the officers came, and found them not in et miserunt in carcerem ut adducerentur 5:22 cum uenis-

the prison, they returned and told,

sent autem ministri et aperto carcere non inuenissent illos

5:23 Saying, The prison truly found we shut with all reuersi nuntiauerunt 5:23 dicentes carcerem quidem inue-

safety, and the keepers standing without before the doors: but when we had opened, we found no man within.

nimus clausum cum omni diligentia et custodes stantes ad

5:24 Now when the high priest and the captain of the tem- ianuas aperientes autem neminem intus inuenimus 5:24 ut

ple and the chief priests heard these things, they doubted of them whereunto this would grow.

audierunt autem hos sermones magistratus templi et prin-

5:25 Then came one and told them, saying, Behold, the cipes sacerdotum ambigebant de illis quidnam fieret 5:25

men whom ye put in prison are standing in the temple, and teaching the people.

adueniens autem quidam nuntiauit eis quia ecce uiri quos posuistis in carcere sunt in templo stantes et docentes po-

5:26 Then went the captain with the officers, and brought pulum 5:26 tunc abiit magistratus cum ministris et adduxit

them without violence: for they feared the people, lest they should have been stoned.

illos sine ui timebant enim populum ne lapidarentur 5:27 et

5:27 And when they had brought them, they set them before the council: and the high priest asked them,

cum adduxissent illos statuerunt in concilio et interrogauit

5:28 Saying, Did not we straitly command you that ye eos princeps sacerdotum 5:28 dicens praecipiendo praece-

should not teach in this name? and, behold, ye have filled Jerusalem with your doctrine, and intend to bring

this man’s blood upon us. pimus uobis ne doceretis in nomine isto et ecce replestis Hie-

rusalem doctrina uestra et uultis inducere super nos sangui-

5:29 Then Peter and the other apostles answered and said, nem hominis istius 5:29 respondens autem Petrus et apo-We ought to obey God rather than men.

stoli dixerunt oboedire oportet Deo magis quam hominibus

5:30 The God of our fathers raised up Jesus, whom ye slew 5:30 Deus patrum nostrorum suscitauit Iesum quem uos in-

and hanged on a tree.

5:31 Him hath God exalted with his right hand to be a teremistis suspendentes in ligno 5:31 hunc Deus principem

Prince and a Saviour, for to give repentance to Israel, and forgiveness of sins.

et saluatorem exaltauit dextera sua ad dandam paenitentiam

5:32 And we are his witnesses of these things; and so is Israhel et remissionem peccatorum 5:32 et nos sumus te-

also the Holy Ghost, whom God hath given to them that obey him.

stes horum uerborum et Spiritus Sanctus quem dedit Deus

5:33 When they heard that, they were cut to the heart, and omnibus oboedientibus sibi 5:33 haec cum audissent disse-

took counsel to slay them.

5:34 Then stood there up one in the council, a Pharisee, cabantur et cogitabant interficere illos 5:34 surgens autem

named Gamaliel, a doctor of the law, had in reputation among all the people, and commanded to put the apostles

forth a little space; quidam in concilio Pharisaeus nomine Gamalihel legis doctor honorabilis uniuersae plebi iussit foras ad breue homines

Actus Apostolorum

fieri 5:35 dixitque ad illos uiri israhelitae adtendite uobis 5:35 And said unto them, Ye men of Israel, take heed to

yourselves what ye intend to do as touching these men.

super hominibus istis quid acturi sitis 5:36 ante hos enim 5:36 For before these days rose up Theudas, boasting him-

self to be somebody; to whom a number of men, about four hundred, joined themselves: who was slain; and all, as many as obeyed him, were scattered, and brought to nought.

dies extitit Theodas dicens esse se aliquem cui consensit uirorum numerus circiter quadringentorum qui occisus est et omnes quicumque credebant ei dissipati sunt et redactus est ad nihilum 5:37 post hunc extitit Iudas Galilaeus in diebus 5:37 After this man rose up Judas of Galilee in the days

of the taxing, and drew away much people after him: he also perished; and all, even as many as obeyed him, were

professionis et auertit populum post se et ipse periit et omnes dispersed. quotquot consenserunt ei dispersi sunt 5:38 et nunc itaque 5:38 And now I say unto you, Refrain from these men, and

let them alone: for if this counsel or this work be of men, it will come to nought:

dico uobis discedite ab hominibus istis et sinite illos quoniam si est ex hominibus consilium hoc aut opus dissoluetur 5:39 si uero ex Deo est non poteritis dissoluere eos ne 5:39 But if it be of God, ye cannot overthrow it; lest haply

ye be found even to fight against God.

forte et Deo repugnare inueniamini consenserunt autem illi 5:40 et conuocantes apostolos caesis denuntiauerunt ne lo- 5:40 And to him they agreed: and when they had called

the apostles, and beaten them, they commanded that they should not speak in the name of Jesus, and let them go.

querentur in nomine Iesu et dimiserunt eos 5:41 et illi qui-

5:41 And they departed from the presence of the council, rejoicing that they were counted worthy to suffer shame for his name.

dem ibant gaudentes a conspectu concilii quoniam digni habiti sunt pro nomine Iesu contumeliam pati 5:42 omni au- 5:42 And daily in the temple, and in every house, they

ceased not to teach and preach Jesus Christ.

tem die in templo et circa domos non cessabant docentes et euangelizantes Christum Iesum

6:1 in diebus autem illis crescente numero discipulo- 6:1 And in those days, when the number of the disciples

was multiplied, there arose a murmuring of the Grecians against the Hebrews, because their widows were neglected

rum factus est murmur Graecorum aduersus Hebraeos eo in the daily ministration. quod dispicerentur in ministerio cotidiano uiduae eorum 6:2 conuocantes autem duodecim multitudinem discipulo- 6:2 Then the twelve called the multitude of the disciples

unto them, and said, It is not reason that we should leave the word of God, and serve tables.

rum dixerunt non est aequum nos derelinquere uerbum Dei et ministrare mensis 6:3 considerate ergo fratres uiros ex 6:3 Wherefore, brethren, look ye out among you seven

men of honest report, full of the Holy Ghost and wisdom, whom we may appoint over this business.

uobis boni testimonii septem plenos Spiritu et sapientia quos

constituamus super hoc opus 6:4 nos uero orationi et mi- 6:4 But we will give ourselves continually to prayer, and

to the ministry of the word.

nisterio uerbi instantes erimus 6:5 et placuit sermo coram 6:5 And the saying pleased the whole multitude: and they

chose Stephen, a man full of faith and of the Holy Ghost, and Philip, and Prochorus, and Nicanor, and Timon, and

omni multitudine et elegerunt Stephanum uirum plenum fide Parmenas, and Nicolas a proselyte of Antioch: et Spiritu Sancto et Philippum et Prochorum et Nicanorem et Timonem et Parmenam et Nicolaum aduenam Antiochenum 6:6 hos statuerunt ante conspectum apostolorum et 6:6 Whom they set before the apostles: and when they had

prayed, they laid their hands on them.

orantes inposuerunt eis manus 6:7 et uerbum Dei cresce- 6:7 And the word of God increased; and the number of

the disciples multiplied in Jerusalem greatly; and a great company of the priests were obedient to the faith.

bat et multiplicabatur numerus discipulorum in Hierusalem

Biblia Sacra Vulgata

6:8 And Stephen, full of faith and power, did great won- ualde multa etiam turba sacerdotum oboediebat fidei 6:8

ders and miracles among the people.

Stephanus autem plenus gratia et fortitudine faciebat prod-

6:9 Then there arose certain of the synagogue, which is iiia et signa magna in populo 6:9 surrexerunt autem quidam

called the synagogue of the Libertines, and Cyrenians, and Alexandrians, and of them of Cilicia and of Asia, disput-

ing with Stephen. de synagoga quae appellatur Libertinorum et Cyrenensium

et Alexandrinorum et eorum qui erant a Cilicia et Asia dispu-

6:10 And they were not able to resist the wisdom and the tantes cum Stephano 6:10 et non poterant resistere sapien-

spirit by which he spake.

6:11 Then they suborned men, which said, We have heard tiae et Spiritui quo loquebatur 6:11 tunc submiserunt uiros

him speak blasphemous words against Moses, and against God.

qui dicerent se audisse eum dicentem uerba blasphemiae in

6:12 And they stirred up the people, and the elders, and the Mosen et Deum 6:12 commouerunt itaque plebem et sescribes, and came upon him, and caught him, and brought him to the council,

niores et scribas et concurrentes rapuerunt eum et adduxe-

6:13 And set up false witnesses, which said, This man runt in concilium 6:13 et statuerunt testes falsos dicentes

ceaseth not to speak blasphemous words against this holy place, and the law:

homo iste non cessat loqui uerba aduersus locum sanctum et

6:14 For we have heard him say, that this Jesus of Nazareth legem 6:14 audiuimus enim eum dicentem quoniam Iesus shall destroy this place, and shall change the customs which Moses delivered us.

Nazarenus hic destruet locum istum et mutabit traditiones

6:15 And all that sat in the council, looking stedfastly on quas tradidit nobis Moses 6:15 et intuentes eum omnes qui him, saw his face as it had been the face of an angel.

sedebant in concilio uiderunt faciem eius tamquam faciem angeli

7:1 Then said the high priest, Are these things so? 7:1 dixit autem princeps sacerdotum si haec ita se habent 7:2 And he said, Men, brethren, and fathers, hearken; The 7:2 qui ait uiri fratres et patres audite Deus gloriae apparuit

God of glory appeared unto our father Abraham, when he was in Mesopotamia, before he dwelt in Charran,

patri nostro Abraham cum esset in Mesopotamiam prius-

7:3 And said unto him, Get thee out of thy country, and quam moraretur in Charram 7:3 et dixit ad illum exi de from thy kindred, and come into the land which I shall shew thee.

terra tua et de cognatione tua et ueni in terram quam tibi

7:4 Then came he out of the land of the Chaldaeans, and monstrauero 7:4 tunc exiit de terra Chaldeorum et habitauit

dwelt in Charran: and from thence, when his father was dead, he removed him into this land, wherein ye now

dwell. in Charram et inde postquam mortuus est pater eius transtu-7:5 And he gave him none inheritance in it, no, not so lit illum in terram istam in qua nunc uos habitatis 7:5 et

much as to set his foot on: yet he promised that he would give it to him for a possession, and to his seed after him,

when as yet he had no child. non dedit illi hereditatem in ea nec passum pedis et repromi-

sit dare illi eam in possessionem et semini eius post ipsum

7:6 And God spake on this wise, That his seed should so- cum non haberet filium 7:6 locutus est autem Deus quia erit

journ in a strange land; and that they should bring them into bondage, and entreat them evil four hundred years.

semen eius accola in terra aliena et seruituti eos subicient

7:7 And the nation to whom they shall be in bondage will et male tractabunt eos annis quadringentis 7:7 et gentem

I judge, said God: and after that shall they come forth, and serve me in this place.

cui seruierint iudicabo ego dixit Deus et post haec exibunt

7:8 And he gave him the covenant of circumcision: and et deseruient mihi in loco isto 7:8 et dedit illi testamen-

so Abraham begat Isaac, and circumcised him the eighth day; and Isaac begat Jacob; and Jacob begat the twelve

patriarchs. tum circumcisionis et sic genuit Isaac et circumcidit eum die

Actus Apostolorum

octaua et Isaac Iacob et Iacob duodecim patriarchas 7:9 et 7:9 And the patriarchs, moved with envy, sold Joseph into

Egypt: but God was with him,

patriarchae aemulantes Ioseph uendiderunt in Aegyptum et

erat Deus cum eo 7:10 et eripuit eum ex omnibus tribula- 7:10 And delivered him out of all his afflictions, and gave him favour and wisdom in the sight of Pharaoh king of Egypt; and he made him governor over Egypt and all his

tionibus eius et dedit ei gratiam et sapientiam in conspectu house. Pharaonis regis Aegypti et constituit eum praepositum super Aegyptum et super omnem domum suam 7:11 uenit au- 7:11 Now there came a dearth over all the land of Egypt

and Chanaan, and great affliction: and our fathers found no sustenance.

tem fames in uniuersam Aegyptum et Chanaan et tribulatio magna et non inueniebant cibos patres nostri 7:12 cum au- 7:12 But when Jacob heard that there was corn in Egypt,

he sent out our fathers first.

disset autem Iacob esse frumentum in Aegypto misit patres nostros primum 7:13 et in secundo cognitus est Ioseph a 7:13 And at the second time Joseph was made known to

his brethren; and Joseph’s kindred was made known unto Pharaoh.

fratribus suis et manifestatum est Pharaoni genus eius 7:14

7:14 Then sent Joseph, and called his father Jacob to him, and all his kindred, threescore and fifteen souls.

mittens autem Ioseph accersiuit Iacob patrem suum et omnem cognationem in animabus septuaginta quinque 7:15 et 7:15 So Jacob went down into Egypt, and died, he, and

our fathers,

descendit Iacob in Aegyptum et defunctus est ipse et patres nostri 7:16 et translati sunt in Sychem et positi sunt in 7:16 And were carried over into Sychem, and laid in the

sepulchre that Abraham bought for a sum of money of the sons of Emmor the father of Sychem.

sepulchro quod emit Abraham pretio argenti a filiis Emmor filii Sychem 7:17 cum adpropinquaret autem tempus repro- 7:17 But when the time of the promise drew nigh, which

God had sworn to Abraham, the people grew and multiplied in Egypt,

missionis quam confessus erat Deus Abrahae creuit populus et multiplicatus est in Aegypto 7:18 quoadusque surrexit 7:18 Till another king arose, which knew not Joseph.

rex alius in Aegypto qui non sciebat Ioseph 7:19 hic cir- 7:19 The same dealt subtilly with our kindred, and evil entreated our fathers, so that they cast out their young children, to the end they might not live.

cumueniens genus nostrum adflixit patres ut exponerent infantes suos ne uiuificarentur 7:20 eodem tempore natus est 7:20 In which time Moses was born, and was exceeding

fair, and nourished up in his father’s house three months:

Moses et fuit gratus Deo qui nutritus est tribus mensibus in domo patris sui 7:21 exposito autem illo sustulit eum filia 7:21 And when he was cast out, Pharaoh’s daughter took

him up, and nourished him for her own son.

Pharaonis et enutriuit eum sibi in filium 7:22 et eruditus est 7:22 And Moses was learned in all the wisdom of the

Egyptians, and was mighty in words and in deeds.

Moses omni sapientia Aegyptiorum et erat potens in uerbis et in operibus suis 7:23 cum autem impleretur ei quadra- 7:23 And when he was full forty years old, it came into his

heart to visit his brethren the children of Israel.

ginta annorum tempus ascendit in cor eius ut uisitaret fratres suos filios Israhel 7:24 et cum uidisset quendam iniuriam 7:24 And seeing one of them suffer wrong, he defended

him, and avenged him that was oppressed, and smote the Egyptian:

patientem uindicauit illum et fecit ultionem ei qui iniuriam sustinebat percusso Aegyptio 7:25 existimabat autem intel- 7:25 For he supposed his brethren would have understood

how that God by his hand would deliver them: but they understood not.

legere fratres quoniam Deus per manum ipsius daret salutem illis at illi non intellexerunt 7:26 sequenti uero die apparuit 7:26 And the next day he shewed himself unto them as

they strove, and would have set them at one again, saying, Sirs, ye are brethren; why do ye wrong one to another?

Biblia Sacra Vulgata

illis litigantibus et reconciliabat eos in pacem dicens uiri fra-

7:27 But he that did his neighbour wrong thrust him away, tres estis ut quid nocetis alterutrum 7:27 qui autem iniuriam

saying, Who made thee a ruler and a judge over us?

faciebat proximo reppulit eum dicens quis te constituit prin-

7:28 Wilt thou kill me, as thou diddest the Egyptian yes- cipem et iudicem super nos 7:28 numquid interficere me tu terday?

7:29 Then fled Moses at this saying, and was a stranger in uis quemadmodum interfecisti heri Aegyptium 7:29 fugit

the land of Madian, where he begat two sons.

autem Moses in uerbo isto et factus est aduena in terra Ma-

7:30 And when forty years were expired, there appeared to diam ubi generauit filios duos 7:30 et expletis annis quad-

him in the wilderness of mount Sina an angel of the Lord in a flame of fire in a bush.

raginta apparuit illi in deserto montis Sina angelus in igne

7:31 When Moses saw it, he wondered at the sight: and flammae rubi 7:31 Moses autem uidens admiratus est ui-

as he drew near to behold it, the voice of the LORD came unto him,

sum et accedente illo ut consideraret facta est uox Domini

7:32 Saying, I am the God of thy fathers, the God of Abra- 7:32 ego Deus patrum tuorum Deus Abraham et Deus Isaac ham, and the God of Isaac, and the God of Jacob. Then Moses trembled, and durst not behold.

et Deus Iacob tremefactus autem Moses non audebat con-

7:33 Then said the Lord to him, Put off thy shoes from thy siderare 7:33 dixit autem illi Dominus solue calciamentum

feet: for the place where thou standest is holy ground.

7:34 I have seen, I have seen the affliction of my people pedum tuorum locus enim in quo stas terra sancta est 7:34

which is in Egypt, and I have heard their groaning, and am come down to deliver them. And now come, I will send

thee into Egypt. uidens uidi adflictionem populi mei qui est in Aegypto et

gemitum eorum audiui et descendi liberare eos et nunc ueni

7:35 This Moses whom they refused, saying, Who made et mittam te in Aegyptum 7:35 hunc Mosen quem negaue-

thee a ruler and a judge? the same did God send to be a ruler and a deliverer by the hand of the angel which ap-

peared to him in the bush. runt dicentes quis te constituit principem et iudicem hunc

Deus principem et redeeptorem misit cum manu angeli qui

7:36 He brought them out, after that he had shewed won- apparuit illi in rubo 7:36 hic eduxit illos faciens prodiiia et

ders and signs in the land of Egypt, and in the Red sea, and in the wilderness forty years.

signa in terra Aegypti et in Rubro mari et in deserto annis

7:37 This is that Moses, which said unto the children of quadraginta 7:37 hic est Moses qui dixit filiis Israhel pro-

Israel, A prophet shall the Lord your God raise up unto you of your brethren, like unto me; him shall ye hear.

phetam uobis suscitabit Deus de fratribus uestris tamquam

7:38 This is he, that was in the church in the wilderness me 7:38 hic est qui fuit in ecclesia in solitudine cum angelo

with the angel which spake to him in the mount Sina, and with our fathers: who received the lively oracles to give

unto us: qui loquebatur ei in monte Sina et cum patribus nostris qui 7:39 To whom our fathers would not obey, but thrust him accepit uerba uitae dare nobis 7:39 cui noluerunt oboedire

from them, and in their hearts turned back again into Egypt,

patres nostri sed reppulerunt et auersi sunt cordibus suis in

7:40 Saying unto Aaron, Make us gods to go before us: Aegyptum 7:40 dicentes ad Aaron fac nobis deos qui prae-

for as for this Moses, which brought us out of the land of Egypt, we wot not what is become of him.

cedant nos Moses enim hic qui eduxit nos de terra Aegypti

7:41 And they made a calf in those days, and offered sac- nescimus quid factum sit ei 7:41 et uitulum fecerunt in il-

rifice unto the idol, and rejoiced in the works of their own hands.

lis diebus et obtulerunt hostiam simulacro et laetabantur in

7:42 Then God turned, and gave them up to worship the operibus manuum suarum 7:42 conuertit autem Deus et tra-

host of heaven; as it is written in the book of the prophets, O ye house of Israel, have ye offered to me slain beasts and

sacrifices by the space of forty years in the wilderness? didit eos seruire militiae caeli sicut scriptum est in libro Pro-

Actus Apostolorum

phetarum numquid uictimas aut hostias obtulistis mihi annis quadraginta in deserto domus Israhel 7:43 et suscepistis 7:43 Yea, ye took up the tabernacle of Moloch, and the star

of your god Remphan, figures which ye made to worship them: and I will carry you away beyond Babylon.

tabernaculum Moloch et sidus dei uestri Rempham figuras quas fecistis adorare eas et transferam uos trans Babylonem 7:44 tabernaculum testimonii fuit patribus nostris in deserto 7:44 Our fathers had the tabernacle of witness in the

wilderness, as he had appointed, speaking unto Moses, that he should make it according to the fashion that he had

sicut disposuit loquens ad Mosen ut faceret illud secundum seen.

formam quam uiderat 7:45 quod et induxerunt suscipien- 7:45 Which also our fathers that came after brought in with Jesus into the possession of the Gentiles, whom God drave out before the face of our fathers, unto the days of David;

tes patres nostri cum Iesu in possessionem gentium quas expulit Deus a facie patrum nostrorum usque in diebus Dauid 7:46 qui inuenit gratiam ante Deum et petiit ut inueniret ta- 7:46 Who found favour before God, and desired to find a

tabernacle for the God of Jacob.

bernaculum Deo Iacob 7:47 Salomon autem aedificauit illi 7:47 But Solomon built him an house. domum 7:48 sed non Excelsus in manufactis habitat sicut 7:48 Howbeit the most High dwelleth not in temples made

with hands; as saith the prophet,

propheta dicit 7:49 caelum mihi sedis est terra autem sca- 7:49 Heaven is my throne, and earth is my footstool: what

house will ye build me? saith the Lord: or what is the place of my rest?

billum pedum meorum quam domum aedificabitis mihi dicit Dominus aut quis locus requietionis meae est 7:50 nonne 7:50 Hath not my hand made all these things? manus mea fecit haec omnia 7:51 dura ceruice et incircum- 7:51 Ye stiffnecked and uncircumcised in heart and ears,

ye do always resist the Holy Ghost: as your fathers did, so do ye.

cisi cordibus et auribus uos semper Spiritui Sancto resistitis sicut patres uestri et uos 7:52 quem prophetarum non sunt 7:52 Which of the prophets have not your fathers perse-

cuted? and they have slain them which shewed before of the coming of the Just One; of whom ye have been now

persecuti patres uestri et occiderunt eos qui praenuntiabant the betrayers and murderers: de aduentu Iusti cuius uos nunc prodiiores et homicidae fuistis 7:53 qui accepistis legem in dispositionem angelorum 7:53 Who have received the law by the disposition of an-

gels, and have not kept it.

et non custodistis 7:54 audientes autem haec dissecabantur 7:54 When they heard these things, they were cut to the

heart, and they gnashed on him with their teeth.

cordibus suis et stridebant dentibus in eum 7:55 cum autem 7:55 But he, being full of the Holy Ghost, looked up sted-

fastly into heaven, and saw the glory of God, and Jesus standing on the right hand of God, and said, Behold, I see the heavens opened, and the Son of man standing on the right hand of God.

esset plenus Spiritu Sancto intendens in caelum uidit gloriam Dei et Iesum stantem a dextris Dei et ait ecce uideo caelos apertos et Filium hominis a dextris stantem Dei 7:56 7:56 Then they cried out with a loud voice, and stopped

their ears, and ran upon him with one accord,

exclamantes autem uoce magna continuerunt aures suas et impetum fecerunt unianimiter in eum 7:57 et eicientes eum 7:57 And cast him out of the city, and stoned him: and the

witnesses laid down their clothes at a young man’s feet, whose name was Saul.

extra ciuitatem lapidabant et testes deposuerunt uestimenta sua secus pedes adulescentis qui uocabatur Saulus 7:58 et 7:58 And they stoned Stephen, calling upon God, and say-

ing, Lord Jesus, receive my spirit.

lapidabant Stephanum inuocantem et dicentem Domine Iesu suscipe spiritum meum 7:59 positis autem genibus clama- 7:59 And he kneeled down, and cried with a loud voice,

Lord, lay not this sin to their charge. And when he had said this, he fell asleep.

uit uoce magna Domine ne statuas illis hoc peccatum et cum

Biblia Sacra Vulgata

hoc dixisset obdormiuit Saulus autem erat consentiens neci eius

8:1 And Saul was consenting unto his death. And at 8:1 facta est autem in illa die persecutio magna in eccle-

that time there was a great persecution against the church which was at Jerusalem; and they were all scattered abroad throughout the regions of Judaea and Samaria, except the apostles.

sia quae erat Hierosolymis et omnes dispersi sunt per regio-

8:2 And devout men carried Stephen to his burial, and nes Iudaeae et Samariae praeter apostolos 8:2 curauerunt

made great lamentation over him.

autem Stephanum uiri timorati et fecerunt planctum mag-

8:3 As for Saul, he made havock of the church, entering num super illum 8:3 Saulus uero deuastabat ecclesiam per

into every house, and haling men and women committed them to prison.

domos intrans et trahens uiros ac mulieres tradebat in cu-

8:4 Therefore they that were scattered abroad went every stodiam 8:4 igitur qui dispersi erant pertransiebant euan-

where preaching the word.

8:5 Then Philip went down to the city of Samaria, and gelizantes uerbum 8:5 Philippus autem descendens in ciui-

preached Christ unto them.

8:6 And the people with one accord gave heed unto those tatem Samariae praedicabat illis Christum 8:6 intendebant

things which Philip spake, hearing and seeing the miracles which he did.

autem turbae his quae a Philippo dicebantur unianimiter au-

8:7 For unclean spirits, crying with loud voice, came out dientes et uidentes signa quae faciebat 8:7 multi enim eo-

of many that were possessed with them: and many taken with palsies, and that were lame, were healed.

rum qui habebant spiritus inmundos clamantes uoce magna

8:8 And there was great joy in that city. exiebant multi autem paralytici et claudi curati sunt 8:8 fac-8:9 But there was a certain man, called Simon, which be- tum est ergo magnum gaudium in illa ciuitate 8:9 uir autem

foretime in the same city used sorcery, and bewitched the people of Samaria, giving out that himself was some great

one: quidam nomine Simon qui ante fuerat in ciuitate magus se-

ducens gentem Samariae dicens esse se aliquem magnum

8:10 To whom they all gave heed, from the least to the 8:10 cui auscultabant omnes a minimo usque ad maximum

greatest, saying, This man is the great power of God.

8:11 And to him they had regard, because that of long time dicentes hic est uirtus Dei quae uocatur Magna 8:11 adten-

he had bewitched them with sorceries.

debant autem eum propter quod multo tempore magicis suis

8:12 But when they believed Philip preaching the things dementasset eos 8:12 cum uero credidissent Philippo euan-

concerning the kingdom of God, and the name of Jesus Christ, they were baptized, both men and women.

gelizanti de regno Dei et nomine Iesu Christi baptizabantur

8:13 Then Simon himself believed also: and when he was uiri ac mulieres 8:13 tunc Simon et ipse credidit et cum ba-

baptized, he continued with Philip, and wondered, beholding the miracles and signs which were done.

ptizatus esset adherebat Philippo uidens etiam signa et uirtu-

8:14 Now when the apostles which were at Jerusalem tes maximas fieri stupens admirabatur 8:14 cum autem au-

heard that Samaria had received the word of God, they sent unto them Peter and John:

dissent apostoli qui erant Hierosolymis quia recepit Sama-

8:15 Who, when they were come down, prayed for them, ria uerbum Dei miserunt ad illos Petrum et Iohannem 8:15

that they might receive the Holy Ghost:

qui cum uenissent orauerunt pro ipsis ut acciperent Spiritum

8:16 (For as yet he was fallen upon none of them: only Sanctum 8:16 nondum enim in quemquam illorum uenerat

they were baptized in the name of the Lord Jesus.)

8:17 Then laid they their hands on them, and they received sed baptizati tantum erant in nomine Domini Iesu 8:17 tunc

the Holy Ghost.

inponebant manus super illos et accipiebant Spiritum Sanc-

8:18 And when Simon saw that through laying on of the tum 8:18 cum uidisset autem Simon quia per inpositionem

apostles’ hands the Holy Ghost was given, he offered them money,

Actus Apostolorum

manus apostolorum daretur Spiritus Sanctus obtulit eis pecuniam 8:19 dicens date et mihi hanc potestatem ut cuicum- 8:19 Saying, Give me also this power, that on whomsoever

I lay hands, he may receive the Holy Ghost.

que inposuero manus accipiat Spiritum Sanctum Petrus autem dixit ad eum 8:20 pecunia tua tecum sit in perditionem 8:20 But Peter said unto him, Thy money perish with thee,

because thou hast thought that the gift of God may be purchased with money.

quoniam donum Dei existimasti pecunia possideri 8:21 non

8:21 Thou hast neither part nor lot in this matter: for thy heart is not right in the sight of God.

est tibi pars neque sors in sermone isto cor enim tuum non est rectum coram Deo 8:22 paenitentiam itaque age ab hac 8:22 Repent therefore of this thy wickedness, and pray

God, if perhaps the thought of thine heart may be forgiven thee.

nequitia tua et roga Deum si forte remittatur tibi haec cogitatio cordis tui 8:23 in felle enim amaritudinis et obliga- 8:23 For I perceive that thou art in the gall of bitterness,

and in the bond of iniquity.

tione iniquitatis uideo te esse 8:24 respondens autem Si- 8:24 Then answered Simon, and said, Pray ye to the

LORD for me, that none of these things which ye have spoken come upon me.

mon dixit precamini uos pro me ad Dominum ut nihil ueniat super me horum quae dixistis 8:25 et illi quidem testificati 8:25 And they, when they had testified and preached the

word of the Lord, returned to Jerusalem, and preached the gospel in many villages of the Samaritans.

et locuti uerbum Domini rediibant Hierosolymam et multis regionibus Samaritanorum euangelizabant 8:26 angelus 8:26 And the angel of the Lord spake unto Philip, saying,

Arise, and go toward the south unto the way that goeth down from Jerusalem unto Gaza, which is desert.

autem Domini locutus est ad Philippum dicens surge et uade contra meridianum ad uiam quae descendit ab Hierusalem in Gazam haec est deserta 8:27 et surgens abiit et ecce uir aet- 8:27 And he arose and went: and, behold, a man of

Ethiopia, an eunuch of great authority under Candace queen of the Ethiopians, who had the charge of all her

hiops eunuchus potens Candacis reginae Aethiopum qui erat treasure, and had come to Jerusalem for to worship, super omnes gazas eius uenerat adorare in Hierusalem 8:28 8:28 Was returning, and sitting in his chariot read Esaias

the prophet.

et reuertebatur sedens super currum suum legensque prophetam Esaiam 8:29 dixit autem Spiritus Philippo accede et ad- 8:29 Then the Spirit said unto Philip, Go near, and join

thyself to this chariot.

iunge te ad currum istum 8:30 adcurrens autem Philippus 8:30 And Philip ran thither to him, and heard him read the

prophet Esaias, and said, Understandest thou what thou readest?

audiuit illum legentem Esaiam prophetam et dixit putasne intellegis quae legis 8:31 qui ait et quomodo possum si non 8:31 And he said, How can I, except some man should

guide me? And he desired Philip that he would come up and sit with him.

aliquis ostenderit mihi rogauitque Philippum ut ascenderet et sederet secum 8:32 locus autem scripturae quam lege- 8:32 The place of the scripture which he read was this, He

was led as a sheep to the slaughter; and like a lamb dumb before his shearer, so opened he not his mouth:

bat erat hic tamquam ouis ad occisionem ductus est et sicut agnus coram tondente se sine uoce sic non aperuit os suum 8:33 in humilitate iudicium eius sublatum est generationem 8:33 In his humiliation his judgment was taken away: and

who shall declare his generation? for his life is taken from the earth.

illius quis enarrabit quoniam tollitur de terra uita eius 8:34

8:34 And the eunuch answered Philip, and said, I pray thee, of whom speaketh the prophet this? of himself, or of some other man?

respondens autem eunuchus Philippo dixit obsecro te de quo propheta dicit hoc de se an de alio aliquo 8:35 aperiens au- 8:35 Then Philip opened his mouth, and began at the same

scripture, and preached unto him Jesus.

tem Philippus os suum et incipiens ab scriptura ista euan-8:36 And as they went on their way, they came unto a cer- gelizauit illi Iesum 8:36 et dum irent per uiam uenerunt ad

tain water: and the eunuch said, See, here is water; what doth hinder me to be baptized?

quandam aquam et ait eunuchus ecce aqua quid prohibet me

8:37 And Philip said, If thou believest with all thine heart, baptizari 8:37 8:38 et iussit stare currum et descenderunt thou mayest. And he answered and said, I believe that Jesus Christ is the Son of God.

8:38 And he commanded the chariot to stand still: and they went down both into the water, both Philip and the eunuch; and he baptized him.

uterque in aquam Philippus et eunuchus et baptizauit eum 8:39 cum autem ascendissent de aqua Spiritus Domini rapuit

8:39 And when they were come up out of the water, the Spirit of the Lord caught away Philip, that the eunuch saw

him no more: and he went on his way rejoicing. Philippum et amplius non uidit eum eunuchus ibat enim per

8:40 But Philip was found at Azotus: and passing through uiam suam gaudens 8:40 Philippus autem inuentus est in he preached in all the cities, till he came to Caesarea.

Azoto et pertransiens euangelizabat ciuitatibus cunctis donec ueniret Caesaream

9:1 And Saul, yet breathing out threatenings and slaughter 9:1 Saulus autem adhuc inspirans minarum et caedis in diagainst the disciples of the Lord, went unto the high priest,

9:2 And desired of him letters to Damascus to the syn- scipulos Domini accessit ad principem sacerdotum 9:2 et

agogues, that if he found any of this way, whether they were men or women, he might bring them bound unto

Jerusalem. petiit ab eo epistulas in Damascum ad synagogas ut si quos inuenisset huius uiae uiros ac mulieres uinctos perduceret in

9:3 And as he journeyed, he came near Damascus: and Hierusalem 9:3 et cum iter faceret contigit ut adpropinqua-

suddenly there shined round about him a light from heaven:

ret Damasco et subito circumfulsit eum lux de caelo 9:4 et

9:4 And he fell to the earth, and heard a voice saying unto him, Saul, Saul, why persecutest thou me?

cadens in terram audiuit uocem dicentem sibi Saule Saule

9:5 And he said, Who art thou, Lord? And the Lord said, quid me persequeris 9:5 qui dixit quis es Domine et ille

I am Jesus whom thou persecutest: it is hard for thee to kick against the pricks.

ego sum Iesus quem tu persequeris 9:6 9:7 sed surge et

9:6 And he trembling and astonished said, Lord, what wilt thou have me to do? And the Lord said unto him, Arise, and go into the city, and it shall be told thee what thou must do.

9:7 And the men which journeyed with him stood speechless, hearing a voice, but seeing no man.

ingredere ciuitatem et dicetur tibi quid te oporteat facere uiri autem illi qui comitabantur cum eo stabant stupefacti au-

9:8 And Saul arose from the earth; and when his eyes were dientes quidem uocem neminem autem uidentes 9:8 suropened, he saw no man: but they led him by the hand, and brought him into Damascus.

rexit autem Saulus de terra apertisque oculis nihil uidebat ad

9:9 And he was three days without sight, and neither did manus autem illum trahentes introduxerunt Damascum 9:9

eat nor drink.

et erat tribus diebus non uidens et non manducauit neque

9:10 And there was a certain disciple at Damascus, named bibit 9:10 erat autem quidam discipulus Damasci nomine Ananias; and to him said the Lord in a vision, Ananias. And he said, Behold, I am here, Lord.

Ananias et dixit ad illum in uisu Dominus Anania at ille ait

9:11 And the Lord said unto him, Arise, and go into the ecce ego Domine 9:11 et Dominus ad illum surgens uade street which is called Straight, and enquire in the house of Judas for one called Saul, of Tarsus: for, behold, he

prayeth, in uicum qui uocatur Rectus et quaere in domo Iudae Sau-9:12 And hath seen in a vision a man named Ananias com- lum nomine Tarsensem ecce enim orat 9:12 et uidit uirum

ing in, and putting his hand on him, that he might receive his sight.

Ananiam nomine introeuntem et inponentem sibi manus ut

9:13 Then Ananias answered, Lord, I have heard by many uisum recipiat 9:13 respondit autem Ananias Domine au-

of this man, how much evil he hath done to thy saints at Jerusalem:

diui a multis de uiro hoc quanta mala sanctis tuis fecerit in

9:14 And here he hath authority from the chief priests to Hierusalem 9:14 et hic habet potestatem a principibus sa-

bind all that call on thy name.

Actus Apostolorum

cerdotum alligandi omnes qui inuocant nomen tuum 9:15 9:15 But the Lord said unto him, Go thy way: for he is a

chosen vessel unto me, to bear my name before the Gentiles, and kings, and the children of Israel:

dixit autem ad eum Dominus uade quoniam uas electionis est mihi iste ut portet nomen meum coram gentibus et regibus et filiis Israhel 9:16 ego enim ostendam illi quanta 9:16 For I will shew him how great things he must suffer

for my name’s sake.

oporteat eum pro nomine meo pati 9:17 et abiit Ananias et 9:17 And Ananias went his way, and entered into the

house; and putting his hands on him said, Brother Saul, the Lord, even Jesus, that appeared unto thee in the way as thou camest, hath sent me, that thou mightest receive thy sight, and be filled with the Holy Ghost.

introiuit in domum et inponens ei manus dixit Saule frater Dominus misit me Iesus qui apparuit tibi in uia qua ueniebas ut uideas et implearis Spiritu Sancto 9:18 et confestim 9:18 And immediately there fell from his eyes as it had

been scales: and he received sight forthwith, and arose, and was baptized.

ceciderunt ab oculis eius tamquam squamae et uisum recepit et surgens baptizatus est 9:19 et cum accepisset cibum 9:19 And when he had received meat, he was strength-

ened. Then was Saul certain days with the disciples which were at Damascus.

confortatus est fuit autem cum discipulis qui erant Damasci per dies aliquot 9:20 et continuo in synagogis praedicabat 9:20 And straightway he preached Christ in the syna-

gogues, that he is the Son of God.

Iesum quoniam hic est Filius Dei 9:21 stupebant autem om- 9:21 But all that heard him were amazed, and said; Is not

this he that destroyed them which called on this name in Jerusalem, and came hither for that intent, that he might

nes qui audiebant et dicebant nonne hic est qui expugnabat bring them bound unto the chief priests? in Hierusalem eos qui inuocabant nomen istud et huc ad hoc uenit ut uinctos illos duceret ad principes sacerdotum 9:22 9:22 But Saul increased the more in strength, and con-

founded the Jews which dwelt at Damascus, proving that this is very Christ.

Saulus autem magis conualescebat et confundebat Iudaeos qui habitabant Damasci adfirmans quoniam hic est Chris-

tus 9:23 cum implerentur autem dies multi consilium fe- 9:23 And after that many days were fulfilled, the Jews took

counsel to kill him:

cerunt Iudaei ut eum interficerent 9:24 notae autem factae 9:24 But their laying await was known of Saul. And they

watched the gates day and night to kill him.

sunt Saulo insidiae eorum custodiebant autem et portas die ac nocte ut eum interficerent 9:25 accipientes autem disci- 9:25 Then the disciples took him by night, and let him

down by the wall in a basket.

puli eius nocte per murum dimiserunt eum submittentes in sporta 9:26 cum autem uenisset in Hierusalem temptabat 9:26 And when Saul was come to Jerusalem, he assayed

to join himself to the disciples: but they were all afraid of him, and believed not that he was a disciple.

iungere se discipulis et omnes timebant eum non credentes quia esset discipulus 9:27 Barnabas autem adprehensum il- 9:27 But Barnabas took him, and brought him to the apos-

tles, and declared unto them how he had seen the Lord in the way, and that he had spoken to him, and how he had

lum duxit ad apostolos et narrauit illis quomodo in uia uidis- preached boldly at Damascus in the name of Jesus. set Dominum et quia locutus est ei et quomodo in Damasco fiducialiter egerit in nomine Iesu 9:28 et erat cum illis in- 9:28 And he was with them coming in and going out at

Jerusalem.

trans et exiens in Hierusalem et fiducialiter agens in nomine Domini 9:29 loquebatur quoque et disputabat cum Grae- 9:29 And he spake boldly in the name of the Lord Jesus,

and disputed against the Grecians: but they went about to slay him.

cis illi autem quaerebant occidere eum 9:30 quod cum co-

9:30 Which when the brethren knew, they brought him down to Caesarea, and sent him forth to Tarsus.

gnouissent fratres deduxerunt eum Caesaream et dimiserunt

Biblia Sacra Vulgata

9:31 Then had the churches rest throughout all Judaea and Tarsum 9:31 ecclesia quidem per totam Iudaeam et Gali-Galilee and Samaria, and were edified; and walking in the fear of the Lord, and in the comfort of the Holy Ghost,

were multiplied. laeam et Samariam habebat pacem et aedificabatur ambu-

lans in timore Domini et consolatione Sancti Spiritus reple-

9:32 And it came to pass, as Peter passed throughout all batur 9:32 factum est autem Petrum dum pertransiret uni-

quarters, he came down also to the saints which dwelt at Lydda.

uersos deuenire et ad sanctos qui habitabant Lyddae 9:33

9:33 And there he found a certain man named Aeneas, which had kept his bed eight years, and was sick of the palsy.

inuenit autem ibi hominem quendam nomine Aeneam ab an-

9:34 And Peter said unto him, Aeneas, Jesus Christ nis octo iacentem in grabatto qui erat paralyticus 9:34 et ait

maketh thee whole: arise, and make thy bed. And he arose immediately.

illi Petrus Aeneas sanat te Iesus Christus surge et sterne tibi

9:35 And all that dwelt at Lydda and Saron saw him, and et continuo surrexit 9:35 et uiderunt illum omnes qui ha-

turned to the Lord.

bitabant Lyddae et Saronae qui conuersi sunt ad Dominum

9:36 Now there was at Joppa a certain disciple named 9:36 in Ioppe autem fuit quaedam discipula nomine Tabitas

Tabitha, which by interpretation is called Dorcas: this woman was full of good works and almsdeeds which she

did. quae interpretata dicitur Dorcas haec erat plena operibus bo-9:37 And it came to pass in those days, that she was sick, nis et elemosynis quas faciebat 9:37 factum est autem in

and died: whom when they had washed, they laid her in an upper chamber.

diebus illis ut infirmata moreretur quam cum lauissent po-

9:38 And forasmuch as Lydda was nigh to Joppa, and the suerunt eam in cenaculo 9:38 cum autem prope esset Lydda

disciples had heard that Peter was there, they sent unto him two men, desiring him that he would not delay to come to

them. ab Ioppe discipuli audientes quia Petrus esset in ea miserunt

duos uiros ad eum rogantes ne pigriteris uenire usque ad nos

9:39 Then Peter arose and went with them. When he was 9:39 exsurgens autem Petrus uenit cum illis et cum aduenisset

come, they brought him into the upper chamber: and all the widows stood by him weeping, and shewing the coats and garments which Dorcas made, while she was with them.

duxerunt illum in cenaculum et circumsteterunt illum omnes uiduae flentes et ostendentes tunicas et uestes quas faciebat

9:40 But Peter put them all forth, and kneeled down, and illis Dorcas 9:40 eiectis autem omnibus foras Petrus ponens

prayed; and turning him to the body said, Tabitha, arise. And she opened her eyes: and when she saw Peter, she sat

up. genua orauit et conuersus ad corpus dixit Tabita surge at illa 9:41 And he gave her his hand, and lifted her up, and when aperuit oculos suos et uiso Petro resedit 9:41 dans autem

he had called the saints and widows, presented her alive.

illi manum erexit eam et cum uocasset sanctos et uiduas ad-

9:42 And it was known throughout all Joppa; and many signauit eam uiuam 9:42 notum autem factum est per uni-

believed in the Lord.

9:43 And it came to pass, that he tarried many days in uersam Ioppen et crediderunt multi in Domino 9:43 factum

Joppa with one Simon a tanner.

est autem ut dies multos moraretur in Ioppe apud quendam Simonem coriarium

10:1 There was a certain man in Caesarea called Cor- 10:1 uir autem quidam erat in Caesarea nomine Cornelius

nelius, a centurion of the band called the Italian band,

10:2 A devout man, and one that feared God with all his centurio cohortis quae dicitur Italica 10:2 religiosus et ti-

house, which gave much alms to the people, and prayed to God alway.

mens Deum cum omni domo sua faciens elemosynas multas

10:3 He saw in a vision evidently about the ninth hour of plebi et deprecans Deum semper 10:3 uidit in uisu manithe day an angel of God coming in to him, and saying unto him, Cornelius.

feste quasi hora nona diei angelum Dei introeuntem ad se et

Actus Apostolorum

dicentem sibi Corneli 10:4 at ille intuens eum timore cor- 10:4 And when he looked on him, he was afraid, and said,

What is it, Lord? And he said unto him, Thy prayers and thine alms are come up for a memorial before God.

reptus dixit quid est domine dixit autem illi orationes tuae et elemosynae tuae ascenderunt in memoriam in conspectu Dei 10:5 et nunc mitte uiros in Ioppen et accersi Simonem 10:5 And now send men to Joppa, and call for one Simon,

whose surname is Peter:

quendam qui cognominatur Petrus 10:6 hic hospitatur apud 10:6 He lodgeth with one Simon a tanner, whose house is

by the sea side: he shall tell thee what thou oughtest to do.

Simonem quendam coriarium cuius est domus iuxta mare 10:7 et cum discessisset angelus qui loquebatur illi uocauit 10:7 And when the angel which spake unto Cornelius was

departed, he called two of his household servants, and a devout soldier of them that waited on him continually;

duos domesticos suos et militem metuentem Dominum ex his qui illi parebant 10:8 quibus cum narrasset omnia misit 10:8 And when he had declared all these things unto them,

he sent them to Joppa.

illos in Ioppen 10:9 postera autem die iter illis facientibus 10:9 On the morrow, as they went on their journey, and

drew nigh unto the city, Peter went up upon the housetop to pray about the sixth hour:

et adpropinquantibus ciuitati ascendit Petrus in superiora ut oraret circa horam sextam 10:10 et cum esuriret uoluit gus- 10:10 And he became very hungry, and would have eaten:

but while they made ready, he fell into a trance,

tare parantibus autem eis cecidit super eum mentis excessus 10:11 et uidet caelum apertum et descendens uas quoddam 10:11 And saw heaven opened, and a certain vessel de-

scending upon him, as it had been a great sheet knit at the four corners, and let down to the earth:

uelut linteum magnum quattuor initiis submitti de caelo in terram 10:12 in quo erant omnia quadrupedia et serpentia 10:12 Wherein were all manner of fourfooted beasts of the

earth, and wild beasts, and creeping things, and fowls of the air.

terrae et uolatilia caeli 10:13 et facta est uox ad eum surge

10:13 And there came a voice to him, Rise, Peter; kill, and eat.

Petre et occide et manduca 10:14 ait autem Petrus absit Do- 10:14 But Peter said, Not so, Lord; for I have never eaten

any thing that is common or unclean.

mine quia numquam manducaui omne commune et inmun-

dum 10:15 et uox iterum secundo ad eum quae Deus purifi- 10:15 And the voice spake unto him again the second time,

What God hath cleansed, that call not thou common.

cauit ne tu commune dixeris 10:16 hoc autem factum est per 10:16 This was done thrice: and the vessel was received

up again into heaven.

ter et statim receptum est uas in caelum 10:17 et dum intra 10:17 Now while Peter doubted in himself what this vision

which he had seen should mean, behold, the men which were sent from Cornelius had made enquiry for Simon’s

se haesitaret Petrus quidnam esset uisio quam uidisset ecce house, and stood before the gate, uiri qui missi erant a Cornelio inquirentes domum Simonis adstiterunt ad ianuam 10:18 et cum uocassent interrogabant 10:18 And called, and asked whether Simon, which was

surnamed Peter, were lodged there.

si Simon qui cognominatur Petrus illic haberet hospitium 10:19 Petro autem cogitante de uisione dixit Spiritus ei ecce 10:19 While Peter thought on the vision, the Spirit said

unto him, Behold, three men seek thee.

uiri tres quaerunt te 10:20 surge itaque et descende et uade 10:20 Arise therefore, and get thee down, and go with

them, doubting nothing: for I have sent them.

cum eis nihil dubitans quia ego misi illos 10:21 descen- 10:21 Then Peter went down to the men which were sent

unto him from Cornelius; and said, Behold, I am he whom ye seek: what is the cause wherefore ye are come?

dens autem Petrus ad uiros dixit ecce ego sum quem quaeritis quae causa est propter quam uenistis 10:22 qui dixerunt 10:22 And they said, Cornelius the centurion, a just man,

and one that feareth God, and of good report among all the nation of the Jews, was warned from God by an holy angel

Cornelius centurio uir iustus et timens Deum et testimonium to send for thee into his house, and to hear words of thee. habens ab uniuersa gente Iudaeorum responsum accepit ab

Biblia Sacra Vulgata

angelo sancto accersire te in domum suam et audire uerba

10:23 Then called he them in, and lodged them. And abs te 10:23 introducens igitur eos recepit hospitio sequenti

on the morrow Peter went away with them, and certain brethren from Joppa accompanied him.

autem die surgens profectus est cum eis et quidam ex fratri-

10:24 And the morrow after they entered into Caesarea. bus ab Ioppe comitati sunt eum 10:24 altera autem die in-

And Cornelius waited for them, and he had called together his kinsmen and near friends.

troiuit Caesaream Cornelius uero expectabat illos conuocatis

10:25 And as Peter was coming in, Cornelius met him, and cognatis suis et necessariis amicis 10:25 et factum est cum

fell down at his feet, and worshipped him.

introisset Petrus obuius ei Cornelius et procidens ad pedes

10:26 But Peter took him up, saying, Stand up; I myself adorauit 10:26 Petrus uero leuauit eum dicens surge et ego

also am a man.

10:27 And as he talked with him, he went in, and found ipse homo sum 10:27 et loquens cum illo intrauit et inuenit

many that were come together.

10:28 And he said unto them, Ye know how that it is an multos qui conuenerant 10:28 dixitque ad illos uos scitis

unlawful thing for a man that is a Jew to keep company, or come unto one of another nation; but God hath shewed

me that I should not call any man common or unclean. quomodo abominatum sit uiro iudaeo coniungi aut accedere

ad alienigenam et mihi ostendit Deus neminem communem

10:29 Therefore came I unto you without gainsaying, as aut inmundum dicere hominem 10:29 propter quod sine du-

soon as I was sent for: I ask therefore for what intent ye have sent for me?

bitatione ueni accersitus interrogo ergo quam ob causam ac-

10:30 And Cornelius said, Four days ago I was fasting cersistis me 10:30 et Cornelius ait a nudius quartana die us-

until this hour; and at the ninth hour I prayed in my house, and, behold, a man stood before me in bright clothing,

que in hanc horam orans eram hora nona in domo mea et

10:31 And said, Cornelius, thy prayer is heard, and thine ecce uir stetit ante me in ueste candida et ait 10:31 Corneli

alms are had in remembrance in the sight of God.

exaudita est oratio tua et elemosynae tuae commemoratae

10:32 Send therefore to Joppa, and call hither Simon, sunt in conspectu Dei 10:32 mitte ergo in Ioppen et accersi

whose surname is Peter; he is lodged in the house of one Simon a tanner by the sea side: who, when he cometh,

shall speak unto thee. Simonem qui cognominatur Petrus hic hospitatur in domo 10:33 Immediately therefore I sent to thee; and thou hast Simonis coriarii iuxta mare 10:33 confestim igitur misi ad

well done that thou art come. Now therefore are we all here present before God, to hear all things that are com-

manded thee of God. te et tu bene fecisti ueniendo nunc ergo omnes nos in con-

spectu tuo adsumus audire omnia quaecumque tibi praecepta

10:34 Then Peter opened his mouth, and said, Of a truth I sunt a Domino 10:34 aperiens autem Petrus os dixit in ue-

perceive that God is no respecter of persons:

ritate conperi quoniam non est personarum acceptor Deus

10:35 But in every nation he that feareth him, and worketh 10:35 sed in omni gente qui timet eum et operatur iustitiam

righteousness, is accepted with him.

10:36 The word which God sent unto the children of Is- acceptus est illi 10:36 uerbum misit filiis Israhel adnuntians

rael, preaching peace by Jesus Christ: (he is Lord of all:)

10:37 That word, I say, ye know, which was published pacem per Iesum Christum hic est omnium Dominus 10:37

throughout all Judaea, and began from Galilee, after the baptism which John preached;

uos scitis quod factum est uerbum per uniuersam Iudaeam incipiens enim a Galilaea post baptismum quod praedicauit

10:38 How God anointed Jesus of Nazareth with the Holy Iohannes 10:38 Iesum a Nazareth quomodo unxit eum De-Ghost and with power: who went about doing good, and healing all that were oppressed of the devil; for God was

with him. us Spiritu Sancto et uirtute qui pertransiuit benefaciendo et sanando omnes oppressos a diabolo quoniam Deus erat cum

Actus Apostolorum

illo 10:39 et nos testes sumus omnium quae fecit in regione 10:39 And we are witnesses of all things which he did both

in the land of the Jews, and in Jerusalem; whom they slew and hanged on a tree:

Iudaeorum et Hierusalem quem et occiderunt suspendentes in ligno 10:40 hunc Deus suscitauit tertia die et dedit eum 10:40 Him God raised up the third day, and shewed him

openly;

manifestum fieri 10:41 non omni populo sed testibus prae- 10:41 Not to all the people, but unto witnesses chosen be-

fore God, even to us, who did eat and drink with him after he rose from the dead.

ordinatis a Deo nobis qui manducauimus et bibimus cum illo postquam resurrexit a mortuis 10:42 et praecepit nobis 10:42 And he commanded us to preach unto the people,

and to testify that it is he which was ordained of God to be the Judge of quick and dead.

praedicare populo et testificari quia ipse est qui constitutus

est a Deo iudex uiuorum et mortuorum 10:43 huic omnes 10:43 To him give all the prophets witness, that through his name whosoever believeth in him shall receive remission of sins.

prophetae testimonium perhibent remissionem peccatorum

accipere per nomen eius omnes qui credunt in eum 10:44 10:44 While Peter yet spake these words, the Holy Ghost

fell on all them which heard the word.

adhuc loquente Petro uerba haec cecidit Spiritus Sanctus super omnes qui audiebant uerbum 10:45 et obstipuerunt ex 10:45 And they of the circumcision which believed were

astonished, as many as came with Peter, because that on the Gentiles also was poured out the gift of the Holy

circumcisione fideles qui uenerant cum Petro quia et in na- Ghost. tiones gratia Spiritus Sancti effusa est 10:46 audiebant enim 10:46 For they heard them speak with tongues, and mag-

nify God. Then answered Peter,

illos loquentes linguis et magnificantes Deum 10:47 tunc 10:47 Can any man forbid water, that these should not be

baptized, which have received the Holy Ghost as well as we?

respondit Petrus numquid aquam quis prohibere potest ut non baptizentur hii qui Spiritum Sanctum acceperunt sicut et nos 10:48 et iussit eos in nomine Iesu Christi baptizari 10:48 And he commanded them to be baptized in the name

of the Lord. Then prayed they him to tarry certain days.

tunc rogauerunt eum ut maneret aliquot diebus

11:1 audierunt autem apostoli et fratres qui erant in Iu- 11:1 And the apostles and brethren that were in Judaea

heard that the Gentiles had also received the word of God.

daea quoniam et gentes receperunt uerbum Dei 11:2 cum 11:2 And when Peter was come up to Jerusalem, they that

were of the circumcision contended with him,

ascendisset autem Petrus in Hierosolymam disceptabant aduersus illum qui erant ex circumcisione 11:3 dicentes quare 11:3 Saying, Thou wentest in to men uncircumcised, and

didst eat with them.

introisti ad uiros praeputium habentes et manducasti cum illis 11:4 incipiens autem Petrus exponebat illis ordinem di- 11:4 But Peter rehearsed the matter from the beginning,

and expounded it by order unto them, saying,

cens 11:5 ego eram in ciuitate Ioppe orans et uidi in ex- 11:5 I was in the city of Joppa praying: and in a trance

I saw a vision, A certain vessel descend, as it had been a great sheet, let down from heaven by four corners; and it

cessu mentis uisionem descendens uas quoddam uelut lin- came even to me: teum magnum quattuor initiis submitti de caelo et uenit usque ad me 11:6 in quod intuens considerabam et uidi qua- 11:6 Upon the which when I had fastened mine eyes, I

considered, and saw fourfooted beasts of the earth, and wild beasts, and creeping things, and fowls of the air.

drupedia terrae et bestias et reptilia et uolatilia caeli 11:7

11:7 And I heard a voice saying unto me, Arise, Peter; slay and eat.

audiui autem et uocem dicentem mihi surgens Petre occide et manduca 11:8 dixi autem nequaquam Domine quia com- 11:8 But I said, Not so, Lord: for nothing common or un-

clean hath at any time entered into my mouth.

mune aut inmundum numquam introiuit in os meum 11:9 11:9 But the voice answered me again from heaven, What God hath cleansed, that call not thou common.

Biblia Sacra Vulgata

respondit autem uox secundo de caelo quae Deus munda-

11:10 And this was done three times: and all were drawn uit tu ne commune dixeris 11:10 hoc autem factum est per

up again into heaven. 11:11 And, behold, immediately there were three men al- ter et recepta sunt rursum omnia in caelum 11:11 et ecce ready come unto the house where I was, sent from Caesarea unto me.

confestim tres uiri adstiterunt in domo in qua eram missi a

11:12 And the Spirit bade me go with them, nothing Caesarea ad me 11:12 dixit autem Spiritus mihi ut irem cum

doubting. Moreover these six brethren accompanied me, and we entered into the man’s house:

illis nihil haesitans uenerunt autem mecum et sex fratres isti

11:13 And he shewed us how he had seen an angel in his et ingressi sumus in domum uiri 11:13 narrauit autem nobis

house, which stood and said unto him, Send men to Joppa, and call for Simon, whose surname is Peter;

quomodo uidisset angelum in domo sua stantem et dicentem sibi mitte in Ioppen et accersi Simonem qui cognominatur

11:14 Who shall tell thee words, whereby thou and all thy Petrus 11:14 qui loquetur tibi uerba in quibus saluus eris

house shall be saved.

11:15 And as I began to speak, the Holy Ghost fell on tu et uniuersa domus tua 11:15 cum autem coepissem lo-

them, as on us at the beginning.

qui decidit Spiritus Sanctus super eos sicut et in nos in initio

11:16 Then remembered I the word of the Lord, how that 11:16 recordatus sum autem uerbi Domini sicut dicebat Io-

he said, John indeed baptized with water; but ye shall be baptized with the Holy Ghost.

hannes quidem baptizauit aqua uos autem baptizabimini Spi-

11:17 Forasmuch then as God gave them the like gift as he ritu Sancto 11:17 si ergo eandem gratiam dedit illis Deus

did unto us, who believed on the Lord Jesus Christ; what was I, that I could withstand God?

sicut et nobis qui credidimus in Dominum Iesum Christum

11:18 When they heard these things, they held their peace, ego quis eram qui possem prohibere Deum 11:18 his auditis and glorified God, saying, Then hath God also to the Gentiles granted repentance unto life.

tacuerunt et glorificauerunt Deum dicentes ergo et gentibus

11:19 Now they which were scattered abroad upon the Deus paenitentiam ad uitam dedit 11:19 et illi quidem qui

persecution that arose about Stephen travelled as far as Phenice, and Cyprus, and Antioch, preaching the word to

none but unto the Jews only. dispersi fuerant a tribulatione quae facta fuerat sub Stephano

perambulauerunt usque Foenicen et Cyprum et Antiochiam

11:20 And some of them were men of Cyprus and Cyrene, nemini loquentes uerbum nisi solis Iudaeis 11:20 erant au-

which, when they were come to Antioch, spake unto the Grecians, preaching the LORD Jesus.

tem quidam ex eis uiri cyprii et cyrenei qui cum introissent Antiochiam loquebantur et ad Graecos adnuntiantes Domi-

11:21 And the hand of the Lord was with them: and a great num Iesum 11:21 et erat manus Domini cum eis multusnumber believed, and turned unto the Lord. 11:22 Then tidings of these things came unto the ears of que numerus credentium conuersus est ad Dominum 11:22 the church which was in Jerusalem: and they sent forth Barnabas, that he should go as far as Antioch.

peruenit autem sermo ad aures ecclesiae quae erat Hierosolymis super istis et miserunt Barnaban usque Antiochiam

11:23 Who, when he came, and had seen the grace of God, 11:23 qui cum peruenisset et uidisset gratiam Dei gauisus est

was glad, and exhorted them all, that with purpose of heart they would cleave unto the Lord.

et hortabatur omnes proposito cordis permanere in Domino

11:24 For he was a good man, and full of the Holy Ghost 11:24 quia erat uir bonus et plenus Spiritu Sancto et fide et

and of faith: and much people was added unto the Lord. 11:25 Then departed Barnabas to Tarsus, for to seek Saul: adposita est turba multa Domino 11:25 profectus est au-

tem Tarsum ut quaereret Saulum quem cum inuenisset per-

Actus Apostolorum

duxit Antiochiam 11:26 et annum totum conuersati sunt in 11:26 And when he had found him, he brought him unto

Antioch. And it came to pass, that a whole year they assembled themselves with the church, and taught much people. And the disciples were called Christians first in Antioch.

ecclesia et docuerunt turbam multam ita ut cognominarentur primum Antiochiae discipuli Christiani 11:27 in his au- 11:27 And in these days came prophets from Jerusalem

unto Antioch.

tem diebus superuenerunt ab Hierosolymis prophetae Antiochiam 11:28 et surgens unus ex eis nomine Agabus signifi- 11:28 And there stood up one of them named Agabus, and

signified by the Spirit that there should be great dearth throughout all the world: which came to pass in the days

cabat per Spiritum famem magnam futuram in uniuerso orbe of Claudius Caesar. terrarum quae facta est sub Claudio 11:29 discipuli autem 11:29 Then the disciples, every man according to his abil-

ity, determined to send relief unto the brethren which dwelt in Judaea:

prout quis habebat proposuerunt singuli eorum in ministerium mittere habitantibus in Iudaea fratribus 11:30 quod et 11:30 Which also they did, and sent it to the elders by the

hands of Barnabas and Saul.

fecerunt mittentes ad seniores per manus Barnabae et Sauli

12:1 eodem autem tempore misit Herodes rex manus ut 12:1 Now about that time Herod the king stretched forth

his hands to vex certain of the church.

adfligeret quosdam de ecclesia 12:2 occidit autem Iacobum 12:2 And he killed James the brother of John with the

sword.

fratrem Iohannis gladio 12:3 uidens autem quia placeret 12:3 And because he saw it pleased the Jews, he proceeded

further to take Peter also. (Then were the days of unleavened bread.)

Iudaeis adposuit adprehendere et Petrum erant autem dies azymorum 12:4 quem cum adprehendisset misit in carce- 12:4 And when he had apprehended him, he put him in

prison, and delivered him to four quaternions of soldiers to keep him; intending after Easter to bring him forth to

rem tradens quattuor quaternionibus militum custodire eum the people. uolens post pascha producere eum populo 12:5 et Petrus 12:5 Peter therefore was kept in prison: but prayer was

made without ceasing of the church unto God for him.

quidem seruabatur in carcere oratio autem fiebat sine intermissione ab ecclesia ad Deum pro eo 12:6 cum autem pro- 12:6 And when Herod would have brought him forth,

the same night Peter was sleeping between two soldiers, bound with two chains: and the keepers before the door

ducturus eum esset Herodes in ipsa nocte erat Petrus dor- kept the prison. miens inter duos milites uinctus catenis duabus et custodes ante ostium custodiebant carcerem 12:7 et ecce angelus 12:7 And, behold, the angel of the Lord came upon him,

and a light shined in the prison: and he smote Peter on the side, and raised him up, saying, Arise up quickly. And his

Domini adstitit et lumen refulsit in habitaculo percussoque chains fell off from his hands. latere Petri suscitauit eum dicens surge uelociter et ceciderunt catenae de manibus eius 12:8 dixit autem angelus ad 12:8 And the angel said unto him, Gird thyself, and bind

on thy sandals. And so he did. And he saith unto him, Cast thy garment about thee, and follow me.

eum praecingere et calcia te gallicas tuas et fecit sic et dixit illi circumda tibi uestimentum tuum et sequere me 12:9 et 12:9 And he went out, and followed him; and wist not that

it was true which was done by the angel; but thought he saw a vision.

exiens sequebatur et nesciebat quia uerum est quod fiebat per angelum aestimabat autem se uisum uidere 12:10 trans- 12:10 When they were past the first and the second ward,

they came unto the iron gate that leadeth unto the city; which opened to them of his own accord: and they went out, and passed on through one street; and forthwith the angel departed from him.

euntes autem primam et secundam custodiam uenerunt ad portam ferream quae ducit ad ciuitatem quae ultro aperta est eis et exeuntes processerunt uicum unum et continuo discessit angelus ab eo 12:11 et Petrus ad se reuersus dixit nunc 12:11 And when Peter was come to himself, he said, Now

I know of a surety, that the LORD hath sent his angel, and hath delivered me out of the hand of Herod, and from all the expectation of the people of the Jews.

Biblia Sacra Vulgata

scio uere quia misit Dominus angelum suum et eripuit me de manu Herodis et de omni expectatione plebis Iudaeorum

12:12 And when he had considered the thing, he came to 12:12 consideransque uenit ad domum Mariae matris Iohan-

the house of Mary the mother of John, whose surname was Mark; where many were gathered together praying.

nis qui cognominatus est Marcus ubi erant multi congregati

12:13 And as Peter knocked at the door of the gate, a et orantes 12:13 pulsante autem eo ostium ianuae processit

damsel came to hearken, named Rhoda.

12:14 And when she knew Peter’s voice, she opened not puella ad audiendum nomine Rhode 12:14 et ut cognouit

the gate for gladness, but ran in, and told how Peter stood before the gate.

uocem Petri prae gaudio non aperuit ianuam sed intro cur-

12:15 And they said unto her, Thou art mad. But she con- rens nuntiauit stare Petrum ante ianuam 12:15 at illi dixe-

stantly affirmed that it was even so. Then said they, It is his angel.

runt ad eam insanis illa autem adfirmabat sic se habere illi

12:16 But Peter continued knocking: and when they had autem dicebant angelus eius est 12:16 Petrus autem perse-

opened the door, and saw him, they were astonished.

uerabat pulsans cum autem aperuissent uiderunt eum et ob-

12:17 But he, beckoning unto them with the hand to hold stipuerunt 12:17 annuens autem eis manu ut tacerent enar-

their peace, declared unto them how the Lord had brought him out of the prison. And he said, Go shew these things unto James, and to the brethren. And he departed, and went into another place.

rauit quomodo Dominus eduxisset eum de carcere dixitque nuntiate Iacobo et fratribus haec et egressus abiit in alium

12:18 Now as soon as it was day, there was no small stir locum 12:18 facta autem die erat non parua turbatio inter

among the soldiers, what was become of Peter.

12:19 And when Herod had sought for him, and found him milites quidnam de Petro factum esset 12:19 Herodes au-

not, he examined the keepers, and commanded that they should be put to death. And he went down from Judaea to

Caesarea, and there abode. tem cum requisisset eum et non inuenisset inquisitione facta

de custodibus iussit eos duci descendensque a Iudaea in Cae-

12:20 And Herod was highly displeased with them of Tyre saream ibi commoratus est 12:20 erat autem iratus Tyriis et

and Sidon: but they came with one accord to him, and, having made Blastus the king’s chamberlain their friend, desired peace; because their country was nourished by the king’s country.

Sidoniis at illi unianimes uenerunt ad eum et persuaso Blasto qui erat super cubiculum regis postulabant pacem eo quod

12:21 And upon a set day Herod, arrayed in royal apparel, alerentur regiones eorum ab illo 12:21 statuto autem die

sat upon his throne, and made an oration unto them.

Herodes uestitus ueste regia sedit pro tribunali et contiona-

12:22 And the people gave a shout, saying, It is the voice batur ad eos 12:22 populus autem adclamabat dei uoces et of a god, and not of a man.

12:23 And immediately the angel of the Lord smote him, non hominis 12:23 confestim autem percussit eum angelus

because he gave not God the glory: and he was eaten of worms, and gave up the ghost.

Domini eo quod non dedisset honorem Deo et consumptus a

12:24 But the word of God grew and multiplied. uermibus exspirauit 12:24 uerbum autem Domini crescebat 12:25 And Barnabas and Saul returned from Jerusalem, et multiplicabatur 12:25 Barnabas autem et Saulus reuersi

when they had fulfilled their ministry, and took with them John, whose surname was Mark.

sunt ab Hierosolymis expleto ministerio adsumpto Iohanne qui cognominatus est Marcus

13:1 Now there were in the church that was at Antioch 13:1 erant autem in ecclesia quae erat Antiochiae prophe-

certain prophets and teachers; as Barnabas, and Simeon that was called Niger, and Lucius of Cyrene, and Manaen, which had been brought up with Herod the tetrarch, and Saul.

tae et doctores in quibus Barnabas et Symeon qui uocabatur Niger et Lucius Cyrenensis et Manaen qui erat Herodis

Actus Apostolorum

tetrarchae conlactaneus et Saulus 13:2 ministrantibus au- 13:2 As they ministered to the Lord, and fasted, the Holy

Ghost said, Separate me Barnabas and Saul for the work whereunto I have called them.

tem illis Domino et ieiunantibus dixit Spiritus Sanctus separate mihi Barnaban et Saulum in opus quod adsumpsi eos 13:3 tunc ieiunantes et orantes inponentesque eis manus di- 13:3 And when they had fasted and prayed, and laid their

hands on them, they sent them away.

miserunt illos 13:4 et ipsi quidem missi ab Spiritu Sancto 13:4 So they, being sent forth by the Holy Ghost, departed

unto Seleucia; and from thence they sailed to Cyprus.

abierunt Seleuciam et inde nauigauerunt Cyprum 13:5 et 13:5 And when they were at Salamis, they preached the

word of God in the synagogues of the Jews: and they had also John to their minister.

cum uenissent Salamina praedicabant uerbum Dei in synagogis Iudaeorum habebant autem et Iohannem in ministerio 13:6 et cum perambulassent uniuersam insulam usque 13:6 And when they had gone through the isle unto Pa-

phos, they found a certain sorcerer, a false prophet, a Jew, whose name was Barjesus:

Paphum inuenerunt quendam uirum magum pseudoprophetam Iudaeum cui nomen erat Bariesu 13:7 qui erat cum pro- 13:7 Which was with the deputy of the country, Sergius

Paulus, a prudent man; who called for Barnabas and Saul, and desired to hear the word of God.

consule Sergio Paulo uiro prudente hic accitis Barnaba et Saulo desiderabat audire uerbum Dei 13:8 resistebat autem 13:8 But Elymas the sorcerer (for so is his name by inter-

pretation) withstood them, seeking to turn away the deputy from the faith.

illis Elymas magus sic enim interpretatur nomen eius quaerens auertere proconsulem a fide 13:9 Saulus autem qui et 13:9 Then Saul, (who also is called Paul,) filled with the

Holy Ghost, set his eyes on him.

Paulus repletus Spiritu Sancto intuens in eum 13:10 dixit o 13:10 And said, O full of all subtilty and all mischief, thou

child of the devil, thou enemy of all righteousness, wilt thou not cease to pervert the right ways of the Lord?

plene omni dolo et omni fallacia fili diaboli inimice omnis iustitiae non desinis subuertere uias Domini rectas 13:11 et 13:11 And now, behold, the hand of the Lord is upon thee,

and thou shalt be blind, not seeing the sun for a season. And immediately there fell on him a mist and a darkness;

nunc ecce manus Domini super te et eris caecus non uidens and he went about seeking some to lead him by the hand. solem usque ad tempus et confestim cecidit in eum caligo et tenebrae et circumiens quaerebat qui ei manum daret 13:12 13:12 Then the deputy, when he saw what was done, be-

lieved, being astonished at the doctrine of the Lord.

tunc proconsul cum uidisset factum credidit admirans super doctrinam Domini 13:13 et cum a Papho nauigassent 13:13 Now when Paul and his company loosed from Pa-

phos, they came to Perga in Pamphylia: and John departing from them returned to Jerusalem.

Paulus et qui cum eo uenerunt Pergen Pamphiliae Iohannes autem discedens ab eis reuersus est Hierosolymam 13:14 13:14 But when they departed from Perga, they came to

Antioch in Pisidia, and went into the synagogue on the sabbath day, and sat down.

illi uero pertranseuntes Pergen uenerunt Antiochiam Pisidiae et ingressi synagogam die sabbatorum sederunt 13:15 13:15 And after the reading of the law and the prophets the

rulers of the synagogue sent unto them, saying, Ye men and brethren, if ye have any word of exhortation for the

post lectionem autem legis et prophetarum miserunt princi- people, say on. pes synagogae ad eos dicentes uiri fratres si quis est in uobis sermo exhortationis ad plebem dicite 13:16 surgens autem 13:16 Then Paul stood up, and beckoning with his hand

said, Men of Israel, and ye that fear God, give audience.

Paulus et manu silentium indicens ait uiri israhelitae et qui timetis Deum audite 13:17 Deus plebis Israhel elegit patres 13:17 The God of this people of Israel chose our fathers,

and exalted the people when they dwelt as strangers in the land of Egypt, and with an high arm brought he them out

nostros et plebem exaltauit cum essent incolae in terra Ae- of it.

Biblia Sacra Vulgata

13:18 And about the time of forty years suffered he their gypti et in brachio excelso eduxit eos ex ea 13:18 et per qua-

manners in the wilderness.

draginta annorum tempus mores eorum sustinuit in deserto

13:19 And when he had destroyed seven nations in the 13:19 et destruens gentes septem in terra Chanaan sorte di-

land of Chanaan, he divided their land to them by lot.

13:20 And after that he gave unto them judges about the stribuit eis terram eorum 13:20 quasi post quadringentos et

space of four hundred and fifty years, until Samuel the prophet.

quinquaginta annos et post haec dedit iudices usque ad Sa-

13:21 And afterward they desired a king: and God gave muhel prophetam 13:21 et exinde postulauerunt regem et

unto them Saul the son of Cis, a man of the tribe of Benjamin, by the space of forty years.

dedit illis Deus Saul filium Cis uirum de tribu Beniamin an-

13:22 And when he had removed him, he raised up unto nis quadraginta 13:22 et amoto illo suscitauit illis Dauid re-

them David to be their king; to whom also he gave their testimony, and said, I have found David the son of Jesse, a

man after mine own heart, which shall fulfil all my will. gem cui et testimonium perhibens dixit inueni Dauid filium

Iesse uirum secundum cor meum qui faciet omnes uolun-

13:23 Of this man’s seed hath God according to his tates meas 13:23 huius Deus ex semine secundum promis-

promise raised unto Israel a Saviour, Jesus:

13:24 When John had first preached before his coming the sionem eduxit Israhel saluatorem Iesum 13:24 praedicante

baptism of repentance to all the people of Israel.

Iohanne ante faciem aduentus eius baptismum paenitentiae

13:25 And as John fulfilled his course, he said, Whom omni populo Israhel 13:25 cum impleret autem Iohannes

think ye that I am? I am not he. But, behold, there cometh one after me, whose shoes of his feet I am not worthy to

loose. cursum suum dicebat quem me arbitramini esse non sum ego

sed ecce uenit post me cuius non sum dignus calciamenta

13:26 Men and brethren, children of the stock of Abraham, pedum soluere 13:26 uiri fratres filii generis Abraham et and whosoever among you feareth God, to you is the word of this salvation sent.

qui in uobis timent Deum uobis uerbum salutis huius mis-

13:27 For they that dwell at Jerusalem, and their rulers, be- sum est 13:27 qui enim habitabant Hierusalem et principes cause they knew him not, nor yet the voices of the prophets which are read every sabbath day, they have fulfilled them

in condemning him. eius hunc ignorantes et uoces prophetarum quae per omne 13:28 And though they found no cause of death in him, sabbatum leguntur iudicantes impleuerunt 13:28 et nullam

yet desired they Pilate that he should be slain.

causam mortis inuenientes in eum petierunt a Pilato ut inter-

13:29 And when they had fulfilled all that was written of ficerent eum 13:29 cumque consummassent omnia quae de him, they took him down from the tree, and laid him in a sepulchre.

eo scripta erant deponentes eum de ligno posuerunt in monu-

13:30 But God raised him from the dead: mento 13:30 Deus uero suscitauit eum a mortuis qui uisus 13:31 And he was seen many days of them which came up est per dies multos his 13:31 qui simul ascenderant cum eo

with him from Galilee to Jerusalem, who are his witnesses unto the people.

de Galilaea in Hierusalem qui usque nunc sunt testes eius

13:32 And we declare unto you glad tidings, how that the ad plebem 13:32 et nos uobis adnuntiamus ea quae ad patpromise which was made unto the fathers,

13:33 God hath fulfilled the same unto us their children, res nostros repromissio facta est 13:33 quoniam hanc Deus

in that he hath raised up Jesus again; as it is also written in the second psalm, Thou art my Son, this day have I

begotten thee. adimpleuit filiis nostris resuscitans Iesum sicut et in psalmo

secundo scriptum est Filius meus es tu ego hodie genui te

13:34 And as concerning that he raised him up from the 13:34 quod autem suscitauerit eum a mortuis amplius iam

dead, now no more to return to corruption, he said on this wise, I will give you the sure mercies of David.

non reuersurum in corruptionem ita dixit quia dabo uobis

Actus Apostolorum

sancta Dauid fidelia 13:35 ideoque et alias dicit non dabis 13:35 Wherefore he saith also in another psalm, Thou

shalt not suffer thine Holy One to see corruption.

Sanctum tuum uidere corruptionem 13:36 Dauid enim sua 13:36 For David, after he had served his own generation

by the will of God, fell on sleep, and was laid unto his fathers, and saw corruption:

generatione cum administrasset uoluntati Dei dormiuit et adpositus est ad patres suos et uidit corruptionem 13:37 quem 13:37 But he, whom God raised again, saw no corruption. uero Deus suscitauit non uidit corruptionem 13:38 notum 13:38 Be it known unto you therefore, men and brethren,

that through this man is preached unto you the forgiveness of sins:

igitur sit uobis uiri fratres quia per hunc uobis remissio peccatorum adnuntiatur ab omnibus quibus non potuistis in lege Mosi iustificari 13:39 in hoc omnis qui credit iustificatur 13:39 And by him all that believe are justified from all

things, from which ye could not be justified by the law of Moses.

13:40 uidete ergo ne superueniat quod dictum est in prophetis

13:40 Beware therefore, lest that come upon you, which is spoken of in the prophets;

13:41 uidete contemptores et admiramini et disperdimini quia 13:41 Behold, ye despisers, and wonder, and perish: for

I work a work in your days, a work which ye shall in no wise believe, though a man declare it unto you.

opus operor ego in diebus uestris opus quod non credetis si quis enarrauerit uobis 13:42 exeuntibus autem illis roga- 13:42 And when the Jews were gone out of the synagogue,

the Gentiles besought that these words might be preached to them the next sabbath.

bant ut sequenti sabbato loquerentur sibi uerba haec 13:43

13:43 Now when the congregation was broken up, many of the Jews and religious proselytes followed Paul and Barnabas: who, speaking to them, persuaded them to continue in the grace of God.

cumque dimissa esset synagoga secuti sunt multi Iudaeorum et colentium aduenarum Paulum et Barnaban qui loquentes suadebant eis ut permanerent in gratia Dei 13:44 sequenti 13:44 And the next sabbath day came almost the whole

city together to hear the word of God.

uero sabbato paene uniuersa ciuitas conuenit audire uerbum Domini 13:45 uidentes autem turbas Iudaei repleti sunt zelo 13:45 But when the Jews saw the multitudes, they were

filled with envy, and spake against those things which were spoken by Paul, contradicting and blaspheming.

et contradicebant his quae a Paulo dicebantur blasphemantes 13:46 tunc constanter Paulus et Barnabas dixerunt uobis 13:46 Then Paul and Barnabas waxed bold, and said, It

was necessary that the word of God should first have been spoken to you: but seeing ye put it from you, and judge yourselves unworthy of everlasting life, lo, we turn to the Gentiles.

oportebat primum loqui uerbum Dei sed quoniam repellitis illud et indignos uos iudicastis aeternae uitae ecce conuertimur ad gentes 13:47 sic enim praecepit nobis Dominus po- 13:47 For so hath the Lord commanded us, saying, I have

set thee to be a light of the Gentiles, that thou shouldest be for salvation unto the ends of the earth.

sui te in lumen gentibus ut sis in salutem usque ad extremum terrae 13:48 audientes autem gentes gauisae sunt et glorifi- 13:48 And when the Gentiles heard this, they were glad,

and glorified the word of the Lord: and as many as were ordained to eternal life believed.

cabant uerbum Domini et crediderunt quotquot erant praeordinati ad uitam aeternam 13:49 disseminabatur autem uer- 13:49 And the word of the Lord was published throughout

all the region.

bum Domini per uniuersam regionem 13:50 Iudaei autem 13:50 But the Jews stirred up the devout and honourable women, and the chief men of the city, and raised persecution against Paul and Barnabas, and expelled them out of

concitauerunt religiosas mulieres et honestas et primos ciui- their coasts. tatis et excitauerunt persecutionem in Paulum et Barnaban et eiecerunt eos de finibus suis 13:51 at illi excusso puluere 13:51 But they shook off the dust of their feet against

them, and came unto Iconium.

pedum in eos uenerunt Iconium 13:52 discipuli quoque re- 13:52 And the disciples were filled with joy, and with the

Holy Ghost.

plebantur gaudio et Spiritu Sancto

Biblia Sacra Vulgata

14:1 And it came to pass in Iconium, that they went both 14:1 factum est autem Iconii ut simul introirent synago-

together into the synagogue of the Jews, and so spake, that a great multitude both of the Jews and also of the Greeks

believed. gam Iudaeorum et loquerentur ita ut crederet Iudaeorum et

14:2 But the unbelieving Jews stirred up the Gentiles, and Graecorum copiosa multitudo 14:2 qui uero increduli fuemade their minds evil affected against the brethren.

runt Iudaei suscitauerunt et ad iracundiam concitauerunt ani-

14:3 Long time therefore abode they speaking boldly in mas gentium aduersus fratres 14:3 multo igitur tempore de-

the Lord, which gave testimony unto the word of his grace, and granted signs and wonders to be done by their hands.

morati sunt fiducialiter agentes in Domino testimonium perhibente uerbo gratiae suae dante signa et prodiiia fieri per

14:4 But the multitude of the city was divided: and part manus eorum 14:4 diuisa est autem multitudo ciuitatis et

held with the Jews, and part with the apostles.

quidam quidem erant cum Iudaeis quidam uero cum apos-

14:5 And when there was an assault made both of the Gen- tolis 14:5 cum autem factus esset impetus gentilium et Iu-

tiles, and also of the Jews with their rulers, to use them despitefully, and to stone them,

daeorum cum principibus suis ut contumeliis adficerent et

14:6 They were ware of it, and fled unto Lystra and Derbe, lapidarent eos 14:6 intellegentes confugerunt ad ciuitates

cities of Lycaonia, and unto the region that lieth round about: And there they preached the gospel.

Lycaoniae Lystram et Derben et uniuersam in circuitu regio-

14:7 And there sat a certain man at Lystra, impotent in his nem et ibi euangelizantes erant 14:7 et quidam uir in Lystris

feet, being a cripple from his mother’s womb, who never had walked:

infirmus pedibus sedebat claudus ex utero matris suae qui

14:8 The same heard Paul speak: who stedfastly beholding numquam ambulauerat 14:8 hic audiuit Paulum loquentem him, and perceiving that he had faith to be healed,

qui intuitus eum et uidens quia haberet fidem ut saluus fieret

14:9 Said with a loud voice, Stand upright on thy feet. And 14:9 dixit magna uoce surge super pedes tuos rectus et exili-

he leaped and walked.

14:10 And when the people saw what Paul had done, they uit et ambulabat 14:10 turbae autem cum uidissent quod fe-

lifted up their voices, saying in the speech of Lycaonia, The gods are come down to us in the likeness of men.

cerat Paulus leuauerunt uocem suam lycaonice dicentes dii

14:11 And they called Barnabas, Jupiter; and Paul, Mer- similes facti hominibus descenderunt ad nos 14:11 et uo-

curius, because he was the chief speaker.

cabant Barnaban Iouem Paulum uero Mercurium quoniam

14:12 Then the priest of Jupiter, which was before their ipse erat dux uerbi 14:12 sacerdos quoque Iouis qui erat city, brought oxen and garlands unto the gates, and would have done sacrifice with the people.

ante ciuitatem tauros et coronas ante ianuas adferens cum

14:13 Which when the apostles, Barnabas and Paul, heard populis uolebat sacrificare 14:13 quod ubi audierunt apos-

of, they rent their clothes, and ran in among the people, crying out,

toli Barnabas et Paulus conscissis tunicis suis exilierunt in

14:14 And saying, Sirs, why do ye these things? We also turbas clamantes 14:14 et dicentes uiri quid haec facitis et

are men of like passions with you, and preach unto you that ye should turn from these vanities unto the living God, which made heaven, and earth, and the sea, and all things that are therein:

nos mortales sumus similes uobis homines adnuntiantes uobis ab his uanis conuerti ad Deum uiuum qui fecit caelum et

14:15 Who in times past suffered all nations to walk in terram et mare et omnia quae in eis sunt 14:15 qui in prae-

their own ways.

teritis generationibus dimisit omnes gentes ingredi in uias

14:16 Nevertheless he left not himself without witness, in suas 14:16 et quidem non sine testimonio semet ipsum rethat he did good, and gave us rain from heaven, and fruitful seasons, filling our hearts with food and gladness.

liquit benefaciens de caelo dans pluuias et tempora fructi-

Actus Apostolorum

fera implens cibo et laetitia corda uestra 14:17 et haec di- 14:17 And with these sayings scarce restrained they the

people, that they had not done sacrifice unto them.

centes uix sedauerunt turbas ne sibi immolarent 14:18 su- 14:18 And there came thither certain Jews from Anti-

och and Iconium, who persuaded the people, and having stoned Paul, drew him out of the city, supposing he had

peruenerunt autem quidam ab Antiochia et Iconio Iudaei et been dead. persuasis turbis lapidantesque Paulum traxerunt extra ciuitatem aestimantes eum mortuum esse 14:19 circumdanti- 14:19 Howbeit, as the disciples stood round about him,

he rose up, and came into the city: and the next day he departed with Barnabas to Derbe.

bus autem eum discipulis surgens intrauit ciuitatem et postera die profectus est cum Barnaba in Derben 14:20 cumque 14:20 And when they had preached the gospel to that city,

and had taught many, they returned again to Lystra, and to Iconium, and Antioch,

euangelizassent ciuitati illi et docuissent multos reuersi sunt

Lystram et Iconium et Antiochiam 14:21 confirmantes ani- 14:21 Confirming the souls of the disciples, and exhorting them to continue in the faith, and that we must through much tribulation enter into the kingdom of God.

mas discipulorum exhortantes ut permanerent in fide et quoniam per multas tribulationes oportet nos intrare in regnum Dei 14:22 et cum constituissent illis per singulas ecclesias 14:22 And when they had ordained them elders in ev-

ery church, and had prayed with fasting, they commended them to the Lord, on whom they believed.

presbyteros et orassent cum ieiunationibus commendauerunt eos Domino in quem crediderunt 14:23 transeuntesque Pi- 14:23 And after they had passed throughout Pisidia, they

came to Pamphylia.

sidiam uenerunt Pamphiliam 14:24 et loquentes in Pergen 14:24 And when they had preached the word in Perga, they

went down into Attalia:

uerbum Domini descenderunt in Attaliam 14:25 et inde na- 14:25 And thence sailed to Antioch, from whence they

had been recommended to the grace of God for the work which they fulfilled.

uigauerunt Antiochiam unde erant traditi gratiae Dei in opus quod conpleuerunt 14:26 cum autem uenissent et congre- 14:26 And when they were come, and had gathered the

church together, they rehearsed all that God had done with them, and how he had opened the door of faith unto the

gassent ecclesiam rettulerunt quanta fecisset Deus cum illis Gentiles. quia aperuisset gentibus ostium fidei 14:27 morati sunt au- 14:27 And there they abode long time with the disciples. tem tempus non modicum cum discipulis

15:1 et quidam descendentes de Iudaea docebant fratres 15:1 And certain men which came down from Judaea

taught the brethren, and said, Except ye be circumcised after the manner of Moses, ye cannot be saved.

quia nisi circumcidamini secundum morem Mosi non potestis salui fieri 15:2 facta ergo seditione non minima Paulo et 15:2 When therefore Paul and Barnabas had no small dis-

sension and disputation with them, they determined that Paul and Barnabas, and certain other of them, should go up to Jerusalem unto the apostles and elders about this question.

Barnabae aduersum illos statuerunt ut ascenderent Paulus et Barnabas et quidam alii ex illis ad apostolos et presbyteros in Hierusalem super hac quaestione 15:3 illi igitur deducti 15:3 And being brought on their way by the church, they

passed through Phenice and Samaria, declaring the conversion of the Gentiles: and they caused great joy unto all

ab ecclesia pertransiebant Foenicen et Samariam narrantes the brethren. conuersionem gentium et faciebant gaudium magnum omnibus fratribus 15:4 cum autem uenissent Hierosolymam su- 15:4 And when they were come to Jerusalem, they were

received of the church, and of the apostles and elders, and they declared all things that God had done with them.

scepti sunt ab ecclesia et ab apostolis et senioribus adnuntiantes quanta Deus fecisset cum illis 15:5 surrexerunt au- 15:5 But there rose up certain of the sect of the Pharisees

which believed, saying, That it was needful to circumcise them, and to command them to keep the law of Moses.

tem quidam de heresi Pharisaeorum qui crediderant dicen-

Biblia Sacra Vulgata

tes quia oportet circumcidi eos praecipere quoque seruare

15:6 And the apostles and elders came together for to con- legem Mosi 15:6 conueneruntque apostoli et seniores uisider of this matter.

15:7 And when there had been much disputing, Peter rose dere de uerbo hoc 15:7 cum autem magna conquisitio fieret

up, and said unto them, Men and brethren, ye know how that a good while ago God made choice among us, that the Gentiles by my mouth should hear the word of the gospel, and believe.

surgens Petrus dixit ad eos uiri fratres uos scitis quoniam ab antiquis diebus in nobis elegit Deus per os meum audire

15:8 And God, which knoweth the hearts, bare them wit- gentes uerbum euangelii et credere 15:8 et qui nouit corda

ness, giving them the Holy Ghost, even as he did unto us;

Deus testimonium perhibuit dans illis Spiritum Sanctum sic-

15:9 And put no difference between us and them, purify- ut et nobis 15:9 et nihil discreuit inter nos et illos fide pu-

ing their hearts by faith.

15:10 Now therefore why tempt ye God, to put a yoke rificans corda eorum 15:10 nunc ergo quid temptatis De-

upon the neck of the disciples, which neither our fathers nor we were able to bear?

um inponere iugum super ceruicem discipulorum quod ne-

15:11 But we believe that through the grace of the LORD que patres nostri neque nos portare potuimus 15:11 sed per

Jesus Christ we shall be saved, even as they.

gratiam Domini Iesu credimus saluari quemadmodum et illi

15:12 Then all the multitude kept silence, and gave audi- 15:12 tacuit autem omnis multitudo et audiebant Barnaban

ence to Barnabas and Paul, declaring what miracles and wonders God had wrought among the Gentiles by them.

et Paulum narrantes quanta fecisset Deus signa et prodiiia

15:13 And after they had held their peace, James an- in gentibus per eos 15:13 et postquam tacuerunt respondit

swered, saying, Men and brethren, hearken unto me:

15:14 Simeon hath declared how God at the first did visit Iacobus dicens uiri fratres audite me 15:14 Simeon narra-

the Gentiles, to take out of them a people for his name.

uit quemadmodum primum Deus uisitauit sumere ex gen-

15:15 And to this agree the words of the prophets; as it is tibus populum nomini suo 15:15 et huic concordant uerba written,

15:16 After this I will return, and will build again the prophetarum sicut scriptum est 15:16 post haec reuertar et

tabernacle of David, which is fallen down; and I will build again the ruins thereof, and I will set it up:

aedificabo tabernaculum Dauid quod decidit et diruta eius

15:17 That the residue of men might seek after the Lord, reaedificabo et erigam illud 15:17 ut requirant ceteri ho-

and all the Gentiles, upon whom my name is called, saith the Lord, who doeth all these things.

minum Dominum et omnes gentes super quas inuocatum

15:18 Known unto God are all his works from the begin- est nomen meum dicit Dominus faciens haec 15:18 notum

ning of the world. 15:19 Wherefore my sentence is, that we trouble not them, a saeculo est Domino opus suum 15:19 propter quod ego which from among the Gentiles are turned to God:

iudico non inquietari eos qui ex gentibus conuertuntur ad

15:20 But that we write unto them, that they abstain from Deum 15:20 sed scribere ad eos ut abstineant se a contapollutions of idols, and from fornication, and from things strangled, and from blood.

minationibus simulacrorum et fornicatione et suffocatis et

15:21 For Moses of old time hath in every city them that sanguine 15:21 Moses enim a temporibus antiquis habet in

preach him, being read in the synagogues every sabbath day.

singulis ciuitatibus qui eum praedicent in synagogis ubi per

15:22 Then pleased it the apostles and elders with the omne sabbatum legitur 15:22 tunc placuit apostolis et se-

whole church, to send chosen men of their own company to Antioch with Paul and Barnabas; namely, Judas sur-

named Barsabas and Silas, chief men among the brethren: nioribus cum omni ecclesia eligere uiros ex eis et mittere

Antiochiam cum Paulo et Barnaba Iudam qui cognominatur

15:23 And they wrote letters by them after this manner; Barsabban et Silam uiros primos in fratribus 15:23 scriben-

The apostles and elders and brethren send greeting unto the brethren which are of the Gentiles in Antioch and Syria and Cilicia.

Actus Apostolorum

tes per manus eorum apostoli et seniores fratres his qui sunt Antiochiae et Syriae et Ciliciae fratribus ex gentibus salutem 15:24 quoniam audiuimus quia quidam ex nobis exeun- 15:24 Forasmuch as we have heard, that certain which

went out from us have troubled you with words, subverting your souls, saying, Ye must be circumcised, and keep

tes turbauerunt uos uerbis euertentes animas uestras quibus the law: to whom we gave no such commandment: non mandauimus 15:25 placuit nobis collectis in unum eli- 15:25 It seemed good unto us, being assembled with one

accord, to send chosen men unto you with our beloved Barnabas and Paul,

gere uiros et mittere ad uos cum carissimis nostris Barnaba et Paulo 15:26 hominibus qui tradiderunt animas suas pro 15:26 Men that have hazarded their lives for the name of

our Lord Jesus Christ.

nomine Domini nostri Iesu Christi 15:27 misimus ergo Iu- 15:27 We have sent therefore Judas and Silas, who shall

also tell you the same things by mouth.

dam et Silam qui et ipsi uobis uerbis referent eadem 15:28 15:28 For it seemed good to the Holy Ghost, and to us,

to lay upon you no greater burden than these necessary things;

uisum est enim Spiritui Sancto et nobis nihil ultra inponere uobis oneris quam haec necessario 15:29 ut abstineatis uos 15:29 That ye abstain from meats offered to idols, and

from blood, and from things strangled, and from fornication: from which if ye keep yourselves, ye shall do well.

ab immolatis simulacrorum et sanguine suffocato et fornica- Fare ye well. tione a quibus custodientes uos bene agetis ualete 15:30 illi 15:30 So when they were dismissed, they came to Anti-

och: and when they had gathered the multitude together, they delivered the epistle:

igitur dimissi descenderunt Antiochiam et congregata multitudine tradiderunt epistulam 15:31 quam cum legissent ga- 15:31 Which when they had read, they rejoiced for the

consolation.

uisi sunt super consolatione 15:32 Iudas autem et Silas et 15:32 And Judas and Silas, being prophets also them-

selves, exhorted the brethren with many words, and confirmed them.

ipsi cum essent prophetae uerbo plurimo consolati sunt fratres et confirmauerunt 15:33 facto autem ibi tempore dimissi 15:33 And after they had tarried there a space, they were

let go in peace from the brethren unto the apostles.

sunt cum pace a fratribus ad eos qui miserant illos 15:34 15:34 Notwithstanding it pleased Silas to abide there still. 15:35 Paulus autem et Barnabas demorabantur Antiochiae 15:35 Paul also and Barnabas continued in Antioch, teach-

ing and preaching the word of the Lord, with many others also.

docentes et euangelizantes cum aliis pluribus uerbum Domini 15:36 post aliquot autem dies dixit ad Barnaban Pau- 15:36 And some days after Paul said unto Barnabas, Let

us go again and visit our brethren in every city where we have preached the word of the LORD, and see how they

lus reuertentes uisitemus fratres per uniuersas ciuitates in do. quibus praedicauimus uerbum Domini quomodo se habeant 15:37 Barnabas autem uolebat secum adsumere et Iohannem 15:37 And Barnabas determined to take with them John,

whose surname was Mark.

qui cognominatur Marcus 15:38 Paulus autem rogabat eum 15:38 But Paul thought not good to take him with them,

who departed from them from Pamphylia, and went not with them to the work.

qui discessisset ab eis a Pamphilia et non isset cum eis in opus non debere recipi eum 15:39 facta est autem dissensio 15:39 And the contention was so sharp between them, that

they departed asunder one from the other: and so Barnabas took Mark, and sailed unto Cyprus;

ita ut discederent ab inuicem et Barnabas adsumpto Marco nauigaret Cyprum 15:40 Paulus uero electo Sila profectus 15:40 And Paul chose Silas, and departed, being recom-

mended by the brethren unto the grace of God.

est traditus gratiae Domini a fratribus 15:41 perambulabat 15:41 And he went through Syria and Cilicia, confirming

the churches.

autem Syriam et Ciliciam confirmans ecclesias

16:1 peruenit autem in Derben et Lystram et ecce disci- 16:1 Then came he to Derbe and Lystra: and, behold, a

certain disciple was there, named Timotheus, the son of a certain woman, which was a Jewess, and believed; but his father was a Greek:

Biblia Sacra Vulgata

pulus quidam erat ibi nomine Timotheus filius mulieris iu-

16:2 Which was well reported of by the brethren that were daeae fidelis patre gentili 16:2 huic testimonium reddebant

at Lystra and Iconium.

16:3 Him would Paul have to go forth with him; and took qui in Lystris erant et Iconii fratres 16:3 hunc uoluit Pau-

and circumcised him because of the Jews which were in those quarters: for they knew all that his father was a

Greek. lus secum proficisci et adsumens circumcidit eum propter

Iudaeos qui erant in illis locis sciebant enim omnes quod

16:4 And as they went through the cities, they delivered pater eius gentilis esset 16:4 cum autem pertransirent ciui-

them the decrees for to keep, that were ordained of the apostles and elders which were at Jerusalem.

tates tradebant eis custodire dogmata quae erant decreta ab

16:5 And so were the churches established in the faith, and apostolis et senioribus qui essent Hierosolymis 16:5 et ec-

increased in number daily.

clesiae quidem confirmabantur fide et abundabant numero

16:6 Now when they had gone throughout Phrygia and the cotidie 16:6 transeuntes autem Frygiam et Galatiae regio-

region of Galatia, and were forbidden of the Holy Ghost to preach the word in Asia,

nem uetati sunt a Sancto Spiritu loqui uerbum in Asia 16:7

16:7 After they were come to Mysia, they assayed to go into Bithynia: but the Spirit suffered them not.

cum uenissent autem in Mysiam temptabant ire Bithyniam et

16:8 And they passing by Mysia came down to Troas. non permisit eos Spiritus Iesu 16:8 cum autem pertransis-16:9 And a vision appeared to Paul in the night; There sent Mysiam descenderunt Troadem 16:9 et uisio per noc-

stood a man of Macedonia, and prayed him, saying, Come over into Macedonia, and help us.

tem Paulo ostensa est uir macedo quidam erat stans et deprecans eum et dicens transiens in Macedoniam adiuua nos

16:10 And after he had seen the vision, immediately we 16:10 ut autem uisum uidit statim quaesiuimus proficisci in

endeavoured to go into Macedonia, assuredly gathering that the Lord had called us for to preach the gospel unto

them. Macedoniam certi facti quia uocasset nos Deus euangelizare 16:11 Therefore loosing from Troas, we came with a eis 16:11 nauigantes autem a Troade recto cursu uenimus

straight course to Samothracia, and the next day to Neapolis;

Samothraciam et sequenti die Neapolim 16:12 et inde Phi-

16:12 And from thence to Philippi, which is the chief city of that part of Macedonia, and a colony: and we were in that city abiding certain days.

lippis quae est prima partis Macedoniae ciuitas colonia era-

16:13 And on the sabbath we went out of the city by a mus autem in hac urbe diebus aliquot conferentes 16:13 die

river side, where prayer was wont to be made; and we sat down, and spake unto the women which resorted thither.

autem sabbatorum egressi sumus foras portam iuxta flumen ubi uidebatur oratio esse et sedentes loquebamur mulieri-

16:14 And a certain woman named Lydia, a seller of pur- bus quae conuenerant 16:14 et quaedam mulier nomine Ly-

ple, of the city of Thyatira, which worshipped God, heard us: whose heart the Lord opened, that she attended unto

the things which were spoken of Paul. dia purpuraria ciuitatis Thyatirenorum colens Deum audiuit

cuius Dominus aperuit cor intendere his quae dicebantur a

16:15 And when she was baptized, and her household, she Paulo 16:15 cum autem baptizata esset et domus eius depre-

besought us, saying, If ye have judged me to be faithful to the Lord, come into my house, and abide there. And she

constrained us. cata est dicens si iudicastis me fidelem Domino esse introite

16:16 And it came to pass, as we went to prayer, a certain in domum meam et manete et coegit nos 16:16 factum est damsel possessed with a spirit of divination met us, which brought her masters much gain by soothsaying:

autem euntibus nobis ad orationem puellam quandam habentem spiritum pythonem obuiare nobis quae quaestum mag-

16:17 The same followed Paul and us, and cried, saying, num praestabat dominis suis diuinando 16:17 haec subse-

These men are the servants of the most high God, which shew unto us the way of salvation.

Actus Apostolorum

cuta Paulum et nos clamabat dicens isti homines serui Dei excelsi sunt qui adnuntiant uobis uiam salutis 16:18 hoc au- 16:18 And this did she many days. But Paul, being

grieved, turned and said to the spirit, I command thee in the name of Jesus Christ to come out of her. And he came

tem faciebat multis diebus dolens autem Paulus et conuersus out the same hour. spiritui dixit praecipio tibi in nomine Iesu Christi exire ab ea et exiit eadem hora 16:19 uidentes autem domini eius quia 16:19 And when her masters saw that the hope of their

gains was gone, they caught Paul and Silas, and drew them into the marketplace unto the rulers,

exiuit spes quaestus eorum adprehendentes Paulum et Silam perduxerunt in forum ad principes 16:20 et offerentes eos 16:20 And brought them to the magistrates, saying, These

men, being Jews, do exceedingly trouble our city,

magistratibus dixerunt hii homines conturbant ciuitatem nostram cum sint Iudaei 16:21 et adnuntiant morem quem non 16:21 And teach customs, which are not lawful for us to

receive, neither to observe, being Romans.

licet nobis suscipere neque facere cum simus Romani 16:22 16:22 And the multitude rose up together against them:

and the magistrates rent off their clothes, and commanded to beat them.

et concurrit plebs aduersus eos et magistratus scissis tunicis eorum iusserunt uirgis caedi 16:23 et cum multas plagas eis 16:23 And when they had laid many stripes upon them,

they cast them into prison, charging the jailor to keep them safely:

inposuissent miserunt eos in carcerem praecipientes custodi ut diligenter custodiret eos 16:24 qui cum tale praeceptum 16:24 Who, having received such a charge, thrust them

into the inner prison, and made their feet fast in the stocks.

accepisset misit eos in interiorem carcerem et pedes eorum strinxit in ligno 16:25 media autem nocte Paulus et Silas 16:25 And at midnight Paul and Silas prayed, and sang

praises unto God: and the prisoners heard them.

adorantes laudabant Deum et audiebant eos qui in custodia erant 16:26 subito uero terraemotus factus est magnus ita ut 16:26 And suddenly there was a great earthquake, so that

the foundations of the prison were shaken: and immediately all the doors were opened, and every one’s bands

mouerentur fundamenta carceris et aperta sunt statim ostia were loosed. omnia et uniuersorum uincula soluta sunt 16:27 experge- 16:27 And the keeper of the prison awaking out of his

sleep, and seeing the prison doors open, he drew out his sword, and would have killed himself, supposing that the

factus autem custos carceris et uidens apertas ianuas carce- prisoners had been fled. ris euaginato gladio uolebat se interficere aestimans fugisse

uinctos 16:28 clamauit autem Paulus magna uoce dicens 16:28 But Paul cried with a loud voice, saying, Do thyself

no harm: for we are all here.

nihil feceris tibi mali uniuersi enim hic sumus 16:29 peti- 16:29 Then he called for a light, and sprang in, and came

trembling, and fell down before Paul and Silas,

toque lumine introgressus est et tremefactus procidit Paulo et Silae 16:30 et producens eos foras ait domini quid me 16:30 And brought them out, and said, Sirs, what must I

do to be saved?

oportet facere ut saluus fiam 16:31 at illi dixerunt crede in 16:31 And they said, Believe on the Lord Jesus Christ, and

thou shalt be saved, and thy house.

Domino Iesu et saluus eris tu et domus tua 16:32 et locuti 16:32 And they spake unto him the word of the Lord, and

to all that were in his house.

sunt ei uerbum Domini cum omnibus qui erant in domo eius 16:33 et tollens eos in illa hora noctis lauit plagas eorum et 16:33 And he took them the same hour of the night, and

washed their stripes; and was baptized, he and all his, straightway.

baptizatus est ipse et omnes eius continuo 16:34 cumque

16:34 And when he had brought them into his house, he set meat before them, and rejoiced, believing in God with all his house.

perduxisset eos in domum suam adposuit eis mensam et lae-

tatus est cum omni domo sua credens Deo 16:35 et cum 16:35 And when it was day, the magistrates sent the serjeants, saying, Let those men go.

Biblia Sacra Vulgata

dies factus esset miserunt magistratus lictores dicentes di-

16:36 And the keeper of the prison told this saying to Paul, mitte homines illos 16:36 nuntiauit autem custos carceris

The magistrates have sent to let you go: now therefore depart, and go in peace.

uerba haec Paulo quia miserunt magistratus ut dimittamini

16:37 But Paul said unto them, They have beaten us nunc igitur exeuntes ite in pace 16:37 Paulus autem dixit eis

openly uncondemned, being Romans, and have cast us into prison; and now do they thrust us out privily? nay

verily; but let them come themselves and fetch us out. caesos nos publice indemnatos homines romanos miserunt

in carcerem et nunc occulte nos eiciunt non ita sed ueniant

16:38 And the serjeants told these words unto the mag- 16:38 et ipsi nos eiciant nuntiauerunt autem magistratibus lic-

istrates: and they feared, when they heard that they were Romans.

tores uerba haec timueruntque audito quod Romani essent

16:39 And they came and besought them, and brought 16:39 et uenientes deprecati sunt eos et educentes rogabant

them out, and desired them to depart out of the city.

16:40 And they went out of the prison, and entered into ut egrederentur urbem 16:40 exeuntes autem de carcere in-

the house of Lydia: and when they had seen the brethren, they comforted them, and departed.

troierunt ad Lydiam et uisis fratribus consolati sunt eos et profecti sunt

17:1 Now when they had passed through Amphipolis and 17:1 cum autem perambulassent Amphipolim et Apollo-

Apollonia, they came to Thessalonica, where was a synagogue of the Jews:

niam uenerunt Thessalonicam ubi erat synagoga Iudaeorum

17:2 And Paul, as his manner was, went in unto them, and 17:2 secundum consuetudinem autem Paulus introiuit ad eos

three sabbath days reasoned with them out of the scriptures,

et per sabbata tria disserebat eis de scripturis 17:3 adape-

17:3 Opening and alleging, that Christ must needs have suffered, and risen again from the dead; and that this Jesus, whom I preach unto you, is Christ.

riens et insinuans quia Christum oportuit pati et resurgere a mortuis et quia hic est Christus Iesus quem ego adnuntio uo-

17:4 And some of them believed, and consorted with Paul bis 17:4 et quidam ex eis crediderunt et adiuncti sunt Paulo

and Silas; and of the devout Greeks a great multitude, and of the chief women not a few.

et Silae et de colentibus gentilibusque multitudo magna et

17:5 But the Jews which believed not, moved with envy, mulieres nobiles non paucae 17:5 zelantes autem Iudaei ad-

took unto them certain lewd fellows of the baser sort, and gathered a company, and set all the city on an uproar, and assaulted the house of Jason, and sought to bring them out to the people.

sumentesque de uulgo uiros quosdam malos et turba facta concitauerunt ciuitatem et adsistentes domui Iasonis quae-

17:6 And when they found them not, they drew Jason and rebant eos producere in populum 17:6 et cum non inuecertain brethren unto the rulers of the city, crying, These that have turned the world upside down are come hither

also; nissent eos trahebant Iasonem et quosdam fratres ad prin-

cipes ciuitatis clamantes quoniam hii qui orbem concitant

17:7 Whom Jason hath received: and these all do contrary et huc uenerunt 17:7 quos suscepit Iason et hii omnes con-

to the decrees of Caesar, saying that there is another king, one Jesus.

tra decreta Caesaris faciunt regem alium dicentes esse Ie-

17:8 And they troubled the people and the rulers of the sum 17:8 concitauerunt autem plebem et principes ciuita-

city, when they heard these things.

17:9 And when they had taken security of Jason, and of tis audientes haec 17:9 et accepto satis ab Iasone et a cete-

the other, they let them go.

17:10 And the brethren immediately sent away Paul and ris dimiserunt eos 17:10 fratres uero confestim per noctem

Silas by night unto Berea: who coming thither went into the synagogue of the Jews.

dimiserunt Paulum et Silam in Beroeam qui cum aduenis-

17:11 These were more noble than those in Thessalonica, sent in synagogam Iudaeorum introierunt 17:11 hii autem

in that they received the word with all readiness of mind, and searched the scriptures daily, whether those things were so.

Actus Apostolorum

erant nobiliores eorum qui sunt Thessalonicae qui susceperunt uerbum cum omni auiditate cotidie scrutantes scripturas si haec ita se haberent 17:12 et multi quidem crediderunt 17:12 Therefore many of them believed; also of hon-

ourable women which were Greeks, and of men, not a few.

ex eis et gentilium mulierum honestarum et uiri non pauci 17:13 cum autem cognouissent in Thessalonica Iudaei quia 17:13 But when the Jews of Thessalonica had knowledge

that the word of God was preached of Paul at Berea, they came thither also, and stirred up the people.

et Beroeae praedicatum est a Paulo uerbum Dei uenerunt et illuc commouentes et turbantes multitudinem 17:14 statim- 17:14 And then immediately the brethren sent away Paul

to go as it were to the sea: but Silas and Timotheus abode there still.

que tunc Paulum dimiserunt fratres ut iret usque ad mare Silas autem et Timotheus remanserunt ibi 17:15 qui autem 17:15 And they that conducted Paul brought him unto

Athens: and receiving a commandment unto Silas and Timotheus for to come to him with all speed, they de-

deducebant Paulum perduxerunt usque Athenas et accepto parted. mandato ab eo ad Silam et Timotheum ut quam celeriter ue-

nirent ad illum profecti sunt 17:16 Paulus autem cum At- 17:16 Now while Paul waited for them at Athens, his spirit was stirred in him, when he saw the city wholly given to idolatry.

henis eos expectaret incitabatur spiritus eius in ipso uidens idolatriae deditam ciuitatem 17:17 disputabat igitur in sy- 17:17 Therefore disputed he in the synagogue with the

Jews, and with the devout persons, and in the market daily with them that met with him.

nagoga cum Iudaeis et colentibus et in foro per omnes dies

ad eos qui aderant 17:18 quidam autem epicurei et stoici 17:18 Then certain philosophers of the Epicureans, and of the Stoicks, encountered him. And some said, What will this babbler say? other some, He seemeth to be a setter forth of strange gods: because he preached unto them Jesus, and the resurrection.

philosophi disserebant cum eo et quidam dicebant quid uult seminiuerbius hic dicere alii uero nouorum daemoniorum uidetur adnuntiator esse quia Iesum et resurrectionem adnuntiabat eis 17:19 et adprehensum eum ad Ariopagum duxe- 17:19 And they took him, and brought him unto Are-

opagus, saying, May we know what this new doctrine, whereof thou speakest, is?

runt dicentes possumus scire quae est haec noua quae a te

dicitur doctrina 17:20 noua enim quaedam infers auribus 17:20 For thou bringest certain strange things to our ears:

we would know therefore what these things mean.

nostris uolumus ergo scire quidnam uelint haec esse 17:21 17:21 (For all the Athenians and strangers which were

there spent their time in nothing else, but either to tell, or to hear some new thing.)

Athenienses autem omnes et aduenae hospites ad nihil aliud uacabant nisi aut dicere aut audire aliquid noui 17:22 17:22 Then Paul stood in the midst of Mars’ hill, and said,

Ye men of Athens, I perceive that in all things ye are too superstitious.

stans autem Paulus in medio Ariopagi ait uiri athenienses per omnia quasi superstitiosiores uos uideo 17:23 praete- 17:23 For as I passed by, and beheld your devotions, I

found an altar with this inscription, TO THE UNKNOWN GOD. Whom therefore ye ignorantly worship, him declare

riens enim et uidens simulacra uestra inueni et aram in qua I unto you. scriptum erat ignoto deo quod ergo ignorantes colitis hoc

ego adnuntio uobis 17:24 Deus qui fecit mundum et om- 17:24 God that made the world and all things therein, seeing that he is Lord of heaven and earth, dwelleth not in temples made with hands;

nia quae in eo sunt hic caeli et terrae cum sit Dominus non in manufactis templis inhabitat 17:25 nec manibus huma- 17:25 Neither is worshipped with men’s hands, as though

he needed any thing, seeing he giveth to all life, and breath, and all things;

nis colitur indigens aliquo cum ipse det omnibus uitam et

Biblia Sacra Vulgata

17:26 And hath made of one blood all nations of men for to inspirationem et omnia 17:26 fecitque ex uno omne genus

dwell on all the face of the earth, and hath determined the times before appointed, and the bounds of their habitation;

hominum inhabitare super uniuersam faciem terrae definiens

17:27 That they should seek the Lord, if haply they might statuta tempora et terminos habitationis eorum 17:27 quae-

feel after him, and find him, though he be not far from every one of us:

rere Deum si forte adtractent eum aut inueniant quamuis non

17:28 For in him we live, and move, and have our being; longe sit ab unoquoque nostrum 17:28 in ipso enim uiui-

as certain also of your own poets have said, For we are also his offspring.

mus et mouemur et sumus sicut et quidam uestrum poetarum

17:29 Forasmuch then as we are the offspring of God, we dixerunt ipsius enim et genus sumus 17:29 genus ergo cum

ought not to think that the Godhead is like unto gold, or silver, or stone, graven by art and man’s device.

simus Dei non debemus aestimare auro aut argento aut lapidi sculpturae artis et cogitationis hominis diuinum esse si-

17:30 And the times of this ignorance God winked at; but mile 17:30 et tempora quidem huius ignorantiae despiciens

now commandeth all men every where to repent:

Deus nunc adnuntiat hominibus ut omnes ubique paeniten-

17:31 Because he hath appointed a day, in the which he tiam agant 17:31 eo quod statuit diem in qua iudicaturus

will judge the world in righteousness by that man whom he hath ordained; whereof he hath given assurance unto all

men, in that he hath raised him from the dead. est orbem in aequitate in uiro in quo statuit fidem praebens

17:32 And when they heard of the resurrection of the dead, omnibus suscitans eum a mortuis 17:32 cum audissent ausome mocked: and others said, We will hear thee again of this matter.

tem resurrectionem mortuorum quidam quidem inridebant

17:33 So Paul departed from among them. quidam uero dixerunt audiemus te de hoc iterum 17:33 sic

17:34 Howbeit certain men clave unto him, and believed: Paulus exiuit de medio eorum 17:34 quidam uero uiri adamong the which was Dionysius the Areopagite, and a woman named Damaris, and others with them.

herentes ei crediderunt in quibus et Dionisius Ariopagita et mulier nomine Damaris et alii cum eis

18:1 After these things Paul departed from Athens, and 18:1 post haec egressus ab Athenis uenit Corinthum 18:2

came to Corinth; 18:2 And found a certain Jew named Aquila, born in Pontus, lately come from Italy, with his wife Priscilla; (because that Claudius had commanded all Jews to depart from Rome:) and came unto them.

et inueniens quendam Iudaeum nomine Aquilam Ponticum genere qui nuper uenerat ab Italia et Priscillam uxorem eius eo quod praecepisset Claudius discedere omnes Iudaeos a

18:3 And because he was of the same craft, he abode with Roma accessit ad eos 18:3 et quia eiusdem erat artis ma-

them, and wrought: for by their occupation they were tentmakers.

nebat apud eos et operabatur erat autem scenofactoriae artis

18:4 And he reasoned in the synagogue every sabbath, and 18:4 18:5 cum uenissent autem de Macedonia Silas et Tipersuaded the Jews and the Greeks. 18:5 And when Silas and Timotheus were come from Macedonia, Paul was pressed in the spirit, and testified to the Jews that Jesus was Christ.

motheus instabat uerbo Paulus testificans Iudaeis esse Chris-

18:6 And when they opposed themselves, and blas- tum Iesum 18:6 contradicentibus autem eis et blaspheman-

phemed, he shook his raiment, and said unto them, Your blood be upon your own heads; I am clean; from hence-

forth I will go unto the Gentiles. tibus excutiens uestimenta dixit ad eos sanguis uester super

18:7 And he departed thence, and entered into a certain caput uestrum mundus ego ex hoc ad gentes uadam 18:7 et man’s house, named Justus, one that worshipped God, whose house joined hard to the synagogue.

migrans inde intrauit in domum cuiusdam nomine Titi Iusti

18:8 And Crispus, the chief ruler of the synagogue, be- colentis Deum cuius domus erat coniuncta synagogae 18:8

lieved on the Lord with all his house; and many of the Corinthians hearing believed, and were baptized.

Crispus autem archisynagogus credidit Domino cum omni

Actus Apostolorum

domo sua et multi Corinthiorum audientes credebant et baptizabantur 18:9 dixit autem Dominus nocte per uisionem 18:9 Then spake the Lord to Paul in the night by a vision,

Be not afraid, but speak, and hold not thy peace:

Paulo noli timere sed loquere et ne taceas 18:10 propter 18:10 For I am with thee, and no man shall set on thee to

hurt thee: for I have much people in this city.

quod ego sum tecum et nemo adponetur tibi ut noceat te quoniam populus est mihi multus in hac ciuitate 18:11 se- 18:11 And he continued there a year and six months,

teaching the word of God among them.

dit autem annum et sex menses docens apud eos uerbum Dei 18:12 Gallione autem proconsule Achaiae insurrexe- 18:12 And when Gallio was the deputy of Achaia, the

Jews made insurrection with one accord against Paul, and brought him to the judgment seat,

runt uno animo Iudaei in Paulum et adduxerunt eum ad tribunal 18:13 dicentes quia contra legem hic persuadet homi- 18:13 Saying, This fellow persuadeth men to worship God

contrary to the law.

nibus colere Deum 18:14 incipiente autem Paulo aperire os 18:14 And when Paul was now about to open his mouth,

Gallio said unto the Jews, If it were a matter of wrong or wicked lewdness, O ye Jews, reason would that I should

dixit Gallio ad Iudaeos si quidem esset iniquum aliquid aut bear with you: facinus pessimum o uiri iudaei recte uos sustinerem 18:15 18:15 But if it be a question of words and names, and of

your law, look ye to it; for I will be no judge of such matters.

si uero quaestiones sunt de uerbo et nominibus et legis uestrae uos ipsi uideritis iudex ego horum nolo esse 18:16 et 18:16 And he drave them from the judgment seat. minauit eos a tribunali 18:17 adprehendentes autem omnes 18:17 Then all the Greeks took Sosthenes, the chief ruler

of the synagogue, and beat him before the judgment seat. And Gallio cared for none of those things.

Sosthenen principem synagogae percutiebant ante tribunal et nihil eorum Gallioni curae erat 18:18 Paulus uero cum 18:18 And Paul after this tarried there yet a good while,

and then took his leave of the brethren, and sailed thence into Syria, and with him Priscilla and Aquila; having shorn

adhuc sustinuisset dies multos fratribus ualefaciens nauiga- his head in Cenchrea: for he had a vow. uit Syriam et cum eo Priscilla et Aquila qui sibi totonderat in Cencris caput habebat enim uotum 18:19 deuenitque 18:19 And he came to Ephesus, and left them there: but

he himself entered into the synagogue, and reasoned with the Jews.

Ephesum et illos ibi reliquit ipse uero ingressus synagogam disputauit cum Iudaeis 18:20 rogantibus autem eis ut amp- 18:20 When they desired him to tarry longer time with

them, he consented not;

liori tempore maneret non consensit 18:21 sed ualefaciens 18:21 But bade them farewell, saying, I must by all means

keep this feast that cometh in Jerusalem: but I will return again unto you, if God will. And he sailed from Ephesus.

et dicens iterum reuertar ad uos Deo uolente profectus est ab Epheso 18:22 et descendens Caesaream ascendit et sa- 18:22 And when he had landed at Caesarea, and gone up,

and saluted the church, he went down to Antioch.

lutauit ecclesiam et descendit Antiochiam 18:23 et facto ibi 18:23 And after he had spent some time there, he departed,

and went over all the country of Galatia and Phrygia in order, strengthening all the disciples.

aliquanto tempore profectus est perambulans ex ordine galaticam regionem et Frygiam confirmans omnes discipulos 18:24 Iudaeus autem quidam Apollo nomine Alexandrinus 18:24 And a certain Jew named Apollos, born at Alexan-

dria, an eloquent man, and mighty in the scriptures, came to Ephesus.

natione uir eloquens deuenit Ephesum potens in scripturis 18:25 hic erat edoctus uiam Domini et feruens spiritu loque- 18:25 This man was instructed in the way of the Lord; and

being fervent in the spirit, he spake and taught diligently the things of the Lord, knowing only the baptism of John.

batur et docebat diligenter ea quae sunt Iesu sciens tantum baptisma Iohannis 18:26 hic ergo coepit fiducialiter agere 18:26 And he began to speak boldly in the synagogue:

whom when Aquila and Priscilla had heard, they took him unto them, and expounded unto him the way of God more perfectly.

Biblia Sacra Vulgata

in synagoga quem cum audissent Priscilla et Aquila adsum-

18:27 And when he was disposed to pass into Achaia, pserunt eum et diligentius exposuerunt ei uiam Dei 18:27

the brethren wrote, exhorting the disciples to receive him: who, when he was come, helped them much which had

believed through grace: cum autem uellet ire Achaiam exhortati fratres scripserunt

discipulis ut susciperent eum qui cum uenisset contulit mul-

18:28 For he mightily convinced the Jews, and that pub- tum his qui crediderant 18:28 uehementer enim Iudaeos re-

lickly, shewing by the scriptures that Jesus was Christ.

uincebat publice ostendens per scripturas esse Christum Iesum

19:1 And it came to pass, that, while Apollos was at 19:1 factum est autem cum Apollo esset Corinthi ut Pau-

Corinth, Paul having passed through the upper coasts came to Ephesus: and finding certain disciples,

lus peragratis superioribus partibus ueniret Ephesum et inue-

19:2 He said unto them, Have ye received the Holy Ghost niret quosdam discipulos 19:2 dixitque ad eos si Spiritum

since ye believed? And they said unto him, We have not so much as heard whether there be any Holy Ghost.

Sanctum accepistis credentes at illi ad eum sed neque si Spi-

19:3 And he said unto them, Unto what then were ye bap- ritus Sanctus est audiuimus 19:3 ille uero ait in quo ergo ba-

tized? And they said, Unto John’s baptism.

19:4 Then said Paul, John verily baptized with the baptism ptizati estis qui dixerunt in Iohannis baptismate 19:4 dixit

of repentance, saying unto the people, that they should believe on him which should come after him, that is, on

Christ Jesus. autem Paulus Iohannes baptizauit baptisma paenitentiae po-

pulum dicens in eum qui uenturus esset post ipsum ut crede-

19:5 When they heard this, they were baptized in the name rent hoc est in Iesum 19:5 his auditis baptizati sunt in no-

of the Lord Jesus.

19:6 And when Paul had laid his hands upon them, the mine Domini Iesu 19:6 et cum inposuisset illis manum Pau-

Holy Ghost came on them; and they spake with tongues, and prophesied.

lus uenit Spiritus Sanctus super eos et loquebantur linguis et

19:7 And all the men were about twelve. prophetabant 19:7 erant autem omnes uiri fere duodecim 19:8 And he went into the synagogue, and spake boldly 19:8 introgressus autem synagogam cum fiducia loquebatur

for the space of three months, disputing and persuading the things concerning the kingdom of God.

per tres menses disputans et suadens de regno Dei 19:9 cum

19:9 But when divers were hardened, and believed not, but spake evil of that way before the multitude, he departed from them, and separated the disciples, disputing daily in the school of one Tyrannus.

autem quidam indurarentur et non crederent maledicentes uiam coram multitudine discedens ab eis segregauit disci-

19:10 And this continued by the space of two years; so pulos cotidie disputans in scola Tyranni 19:10 hoc autem

that all they which dwelt in Asia heard the word of the Lord Jesus, both Jews and Greeks.

factum est per biennium ita ut omnes qui habitabant in Asia

19:11 And God wrought special miracles by the hands of audirent uerbum Domini Iudaei atque gentiles 19:11 uirtu-

Paul:

19:12 So that from his body were brought unto the sick tesque non quaslibet Deus faciebat per manus Pauli 19:12

handkerchiefs or aprons, and the diseases departed from them, and the evil spirits went out of them.

ita ut etiam super languidos deferrentur a corpore eius sudaria uel semicintia et recedebant ab eis languores et spi-

19:13 Then certain of the vagabond Jews, exorcists, took ritus nequam egrediebantur 19:13 temptauerunt autem qui-

upon them to call over them which had evil spirits the name of the LORD Jesus, saying, We adjure you by Je-

sus whom Paul preacheth. dam et de circumeuntibus iudaeis exorcistis inuocare super

eos qui habebant spiritus malos nomen Domini Iesu dicen-

19:14 And there were seven sons of one Sceva, a Jew, and tes adiuro uos per Iesum quem Paulus praedicat 19:14 erant

chief of the priests, which did so.

Actus Apostolorum

autem quidam Sceuae Iudaei principis sacerdotum septem filii qui hoc faciebant 19:15 respondens autem spiritus ne- 19:15 And the evil spirit answered and said, Jesus I know,

and Paul I know; but who are ye?

quam dixit eis Iesum noui et Paulum scio uos autem qui estis 19:16 et insiliens homo in eos in quo erat daemonium pessi- 19:16 And the man in whom the evil spirit was leaped on

them, and overcame them, and prevailed against them, so that they fled out of that house naked and wounded.

mum et dominatus amborum inualuit contra eos ita ut nudi et uulnerati effugerent de domo illa 19:17 hoc autem notum 19:17 And this was known to all the Jews and Greeks also

dwelling at Ephesus; and fear fell on them all, and the name of the Lord Jesus was magnified.

factum est omnibus Iudaeis atque gentilibus qui habitabant Ephesi et cecidit timor super omnes illos et magnificabatur nomen Domini Iesu 19:18 multique credentium ueniebant 19:18 And many that believed came, and confessed, and

shewed their deeds.

confitentes et adnuntiantes actus suos 19:19 multi autem ex 19:19 Many of them also which used curious arts brought

their books together, and burned them before all men: and they counted the price of them, and found it fifty thousand

his qui fuerant curiosa sectati contulerunt libros et conbuse- pieces of silver. runt coram omnibus et conputatis pretiis illorum inuenerunt pecuniam denariorum quinquaginta milium 19:20 ita for- 19:20 So mightily grew the word of God and prevailed. titer uerbum Dei crescebat et confirmabatur 19:21 his au- 19:21 After these things were ended, Paul purposed in

the spirit, when he had passed through Macedonia and Achaia, to go to Jerusalem, saying, After I have been there,

tem expletis posuit Paulus in Spiritu transita Macedonia et I must also see Rome. Achaia ire Hierosolymam dicens quoniam postquam fuero ibi oportet me et Romam uidere 19:22 mittens autem in 19:22 So he sent into Macedonia two of them that min-

istered unto him, Timotheus and Erastus; but he himself stayed in Asia for a season.

Macedoniam duos ex ministrantibus sibi Timotheum et Erastum ipse remansit ad tempus in Asia 19:23 facta est autem 19:23 And the same time there arose no small stir about

that way.

in illo tempore turbatio non minima de uia 19:24 Demet- 19:24 For a certain man named Demetrius, a silversmith, which made silver shrines for Diana, brought no small gain unto the craftsmen;

rius enim quidam nomine argentarius faciens aedes argenteas Dianae praestabat artificibus non modicum quaestum 19:25 quos conuocans et eos qui eiusmodi erant opifices dixit 19:25 Whom he called together with the workmen of like

occupation, and said, Sirs, ye know that by this craft we have our wealth.

uiri scitis quia de hoc artificio adquisitio est nobis 19:26

19:26 Moreover ye see and hear, that not alone at Ephesus, but almost throughout all Asia, this Paul hath persuaded and turned away much people, saying that they be no gods, which are made with hands:

et uidetis et auditis quia non solum Ephesi sed paene totius Asiae Paulus hic suadens auertit multam turbam dicens quo-

niam non sunt dii qui manibus fiunt 19:27 non solum au- 19:27 So that not only this our craft is in danger to be set at nought; but also that the temple of the great goddess Diana should be despised, and her magnificence should be

tem haec periclitabitur nobis pars in redargutionem uenire destroyed, whom all Asia and the world worshippeth. sed et magnae deae Dianae templum in nihilum reputabitur sed et destrui incipiet maiestas eius quam tota Asia et orbis colit 19:28 his auditis repleti sunt ira et exclamauerunt 19:28 And when they heard these sayings, they were full

of wrath, and cried out, saying, Great is Diana of the Ephesians.

dicentes magna Diana Ephesiorum 19:29 et impleta est ci-

19:29 And the whole city was filled with confusion: and having caught Gaius and Aristarchus, men of Macedonia, Paul’s companions in travel, they rushed with one accord into the theatre.

uitas confusione et impetum fecerunt uno animo in theat-

Biblia Sacra Vulgata

rum rapto Gaio et Aristarcho Macedonibus comitibus Pauli

19:30 And when Paul would have entered in unto the peo- 19:30 Paulo autem uolente intrare in populum non permise-

ple, the disciples suffered him not.

19:31 And certain of the chief of Asia, which were his runt discipuli 19:31 quidam autem et de Asiae principibus

friends, sent unto him, desiring him that he would not adventure himself into the theatre.

qui erant amici eius miserunt ad eum rogantes ne se daret in

19:32 Some therefore cried one thing, and some another: theatrum 19:32 alii autem aliud clamabant erat enim eccle-

for the assembly was confused: and the more part knew not wherefore they were come together.

sia confusa et plures nesciebant qua ex causa conuenissent

19:33 And they drew Alexander out of the multitude, the 19:33 de turba autem detraxerunt Alexandrum propellentibus

Jews putting him forward. And Alexander beckoned with the hand, and would have made his defence unto the peo-

ple. eum Iudaeis Alexander ergo manu silentio postulato uole-19:34 But when they knew that he was a Jew, all with one bat rationem reddere populo 19:34 quem ut cognouerunt Iu-

voice about the space of two hours cried out, Great is Diana of the Ephesians.

daeum esse uox facta est una omnium quasi per horas duas

19:35 And when the townclerk had appeased the peo- clamantium magna Diana Ephesiorum 19:35 et cum sedas-

ple, he said, Ye men of Ephesus, what man is there that knoweth not how that the city of the Ephesians is a worshipper of the great goddess Diana, and of the image which fell down from Jupiter?

set scriba turbas dixit uiri ephesii quis enim est hominum qui nesciat Ephesiorum ciuitatem cultricem esse magnae Dia-

19:36 Seeing then that these things cannot be spoken nae Iouisque prolis 19:36 cum ergo his contradici non pos-

against, ye ought to be quiet, and to do nothing rashly.

19:37 For ye have brought hither these men, which are sit oportet uos sedatos esse et nihil temere agere 19:37 ad-

neither robbers of churches, nor yet blasphemers of your goddess.

duxistis enim homines istos neque sacrilegos neque blasphe-

19:38 Wherefore if Demetrius, and the craftsmen which mantes deam uestram 19:38 quod si Demetrius et qui cum are with him, have a matter against any man, the law is open, and there are deputies: let them implead one an-

other. eo sunt artifices habent aduersus aliquem causam conuen-

tus forenses aguntur et pro consulibus sunt accusent inuicem

19:39 But if ye enquire any thing concerning other mat- 19:39 si quid autem alterius rei quaeritis in legitima ecclesia

ters, it shall be determined in a lawful assembly.

19:40 For we are in danger to be called in question for poterit absolui 19:40 nam et periclitamur argui seditionis

this day’s uproar, there being no cause whereby we may give an account of this concourse. And when he had thus

spoken, he dismissed the assembly. hodiernae cum nullus obnoxius sit de quo non possimus reddere rationem concursus istius et cum haec dixisset dimisit ecclesiam

20:1 And after the uproar was ceased, Paul called unto him 20:1 postquam autem cessauit tumultus uocatis Paulus di-

the disciples, and embraced them, and departed for to go into Macedonia.

scipulis et exhortatus eos ualedixit et profectus est ut iret

20:2 And when he had gone over those parts, and had in Macedoniam 20:2 cum autem perambulasset partes illas

given them much exhortation, he came into Greece,

et exhortatus eos fuisset multo sermone uenit ad Graeciam

20:3 And there abode three months. And when the Jews 20:3 ubi cum fecisset menses tres factae sunt illi insidiae a

laid wait for him, as he was about to sail into Syria, he purposed to return through Macedonia.

Iudaeis nauigaturo in Syriam habuitque consilium ut reuer-

20:4 And there accompanied him into Asia Sopater of teretur per Macedoniam 20:4 comitatus est autem eum So-

Berea; and of the Thessalonians, Aristarchus and Secundus; and Gaius of Derbe, and Timotheus; and of Asia, Ty-

chicus and Trophimus. pater Pyrri Beroensis Thessalonicensium uero Aristarchus et Secundus et Gaius Derbeus et Timotheus Asiani uero Tychi-

Actus Apostolorum

cus et Trophimus 20:5 hii cum praecessissent sustinebant 20:5 These going before tarried for us at Troas. nos Troade 20:6 nos uero nauigauimus post dies azymo- 20:6 And we sailed away from Philippi after the days of

unleavened bread, and came unto them to Troas in five days; where we abode seven days.

rum a Philippis et uenimus ad eos Troadem in diebus quinque ubi demorati sumus diebus septem 20:7 in una autem 20:7 And upon the first day of the week, when the dis-

ciples came together to break bread, Paul preached unto them, ready to depart on the morrow; and continued his

sabbati cum conuenissemus ad frangendum panem Paulus speech until midnight. disputabat eis profecturus in crastinum protraxitque sermonem usque in mediam noctem 20:8 erant autem lampades 20:8 And there were many lights in the upper chamber,

where they were gathered together.

copiosae in cenaculo ubi eramus congregati 20:9 sedens au- 20:9 And there sat in a window a certain young man

named Eutychus, being fallen into a deep sleep: and as Paul was long preaching, he sunk down with sleep, and

tem quidam adulescens nomine Eutychus super fenestram fell down from the third loft, and was taken up dead. cum mergeretur somno graui disputante diu Paulo eductus somno cecidit de tertio cenaculo deorsum et sublatus est mortuus 20:10 ad quem cum descendisset Paulus incubuit 20:10 And Paul went down, and fell on him, and embrac-

ing him said, Trouble not yourselves; for his life is in him.

super eum et conplexus dixit nolite turbari anima enim ipsius in eo est 20:11 ascendens autem frangensque panem et 20:11 When he therefore was come up again, and had bro-

ken bread, and eaten, and talked a long while, even till break of day, so he departed.

gustans satisque adlocutus usque in lucem sic profectus est 20:12 adduxerunt autem puerum uiuentem et consolati sunt 20:12 And they brought the young man alive, and were not

a little comforted.

non minime 20:13 nos autem ascendentes nauem enauiga- 20:13 And we went before to ship, and sailed unto Assos,

there intending to take in Paul: for so had he appointed, minding himself to go afoot.

uimus in Asson inde suscepturi Paulum sic enim disposuerat ipse per terram iter facturus 20:14 cum autem conuenisset 20:14 And when he met with us at Assos, we took him in,

and came to Mitylene.

nos in Asson adsumpto eo uenimus Mytilenen 20:15 et in- 20:15 And we sailed thence, and came the next day over

against Chios; and the next day we arrived at Samos, and tarried at Trogyllium; and the next day we came to Mile-

de nauigantes sequenti die uenimus contra Chium et alia ad- tus.

plicuimus Samum et sequenti uenimus Miletum 20:16 pro- 20:16 For Paul had determined to sail by Ephesus, because he would not spend the time in Asia: for he hasted, if it were possible for him, to be at Jerusalem the day of Pen-

posuerat enim Paulus transnauigare Ephesum ne qua mora tecost. illi fieret in Asia festinabat enim si possibile sibi esset ut diem pentecosten faceret Hierosolymis 20:17 a Mileto au- 20:17 And from Miletus he sent to Ephesus, and called the

elders of the church.

tem mittens Ephesum uocauit maiores natu ecclesiae 20:18 20:18 And when they were come to him, he said unto

them, Ye know, from the first day that I came into Asia, after what manner I have been with you at all seasons,

qui cum uenissent ad eum et simul essent dixit eis uos scitis a prima die qua ingressus sum in Asiam qualiter uobiscum per omne tempus fuerim 20:19 seruiens Domino cum omni 20:19 Serving the LORD with all humility of mind, and

with many tears, and temptations, which befell me by the lying in wait of the Jews:

humilitate et lacrimis et temptationibus quae mihi acciderunt ex insidiis Iudaeorum 20:20 quomodo nihil subtraxe- 20:20 And how I kept back nothing that was profitable

unto you, but have shewed you, and have taught you publickly, and from house to house,

rim utilium quo minus adnuntiarem uobis et docerem uos publice et per domos 20:21 testificans Iudaeis atque genti- 20:21 Testifying both to the Jews, and also to the Greeks,

repentance toward God, and faith toward our Lord Jesus Christ.

Biblia Sacra Vulgata

libus in Deum paenitentiam et fidem in Dominum nostrum

20:22 And now, behold, I go bound in the spirit unto Iesum Christum 20:22 et nunc ecce alligatus ego Spiritu

Jerusalem, not knowing the things that shall befall me there:

uado in Hierusalem quae in ea euentura sint mihi ignorans

20:23 Save that the Holy Ghost witnesseth in every city, 20:23 nisi quod Spiritus Sanctus per omnes ciuitates prote-

saying that bonds and afflictions abide me.

statur mihi dicens quoniam uincula et tribulationes me ma-

20:24 But none of these things move me, neither count I nent 20:24 sed nihil horum uereor nec facio animam pre-

my life dear unto myself, so that I might finish my course with joy, and the ministry, which I have received of the Lord Jesus, to testify the gospel of the grace of God. tiosiorem quam me dummodo consummem cursum meum

et ministerium quod accepi a Domino Iesu testificari euan-

20:25 And now, behold, I know that ye all, among whom gelium gratiae Dei 20:25 et nunc ecce ego scio quia amp-

I have gone preaching the kingdom of God, shall see my face no more.

lius non uidebitis faciem meam uos omnes per quos trans-

20:26 Wherefore I take you to record this day, that I am iui praedicans regnum Dei 20:26 quapropter contestor uos

pure from the blood of all men. 20:27 For I have not shunned to declare unto you all the hodierna die quia mundus sum a sanguine omnium 20:27 counsel of God.

non enim subterfugi quo minus adnuntiarem omne consi-

20:28 Take heed therefore unto yourselves, and to all the lium Dei uobis 20:28 adtendite uobis et uniuerso gregi in flock, over the which the Holy Ghost hath made you overseers, to feed the church of God, which he hath purchased

with his own blood. quo uos Spiritus Sanctus posuit episcopos regere ecclesiam 20:29 For I know this, that after my departing shall Dei quam adquisiuit sanguine suo 20:29 ego scio quoniam

grievous wolves enter in among you, not sparing the flock.

intrabunt post discessionem meam lupi graues in uos non

20:30 Also of your own selves shall men arise, speaking parcentes gregi 20:30 et ex uobis ipsis exsurgent uiri lo-

perverse things, to draw away disciples after them.

20:31 Therefore watch, and remember, that by the space quentes peruersa ut abducant discipulos post se 20:31 prop-

of three years I ceased not to warn every one night and day with tears.

ter quod uigilate memoria retinentes quoniam per triennium nocte et die non cessaui cum lacrimis monens unumquem-

20:32 And now, brethren, I commend you to God, and que uestrum 20:32 et nunc commendo uos Deo et uerbo

to the word of his grace, which is able to build you up, and to give you an inheritance among all them which are

sanctified. gratiae ipsius qui potens est aedificare et dare hereditatem 20:33 I have coveted no man’s silver, or gold, or apparel. in sanctificatis omnibus 20:33 argentum aut aurum aut ues-20:34 Yea, ye yourselves know, that these hands have min- tem nullius concupiui 20:34 ipsi scitis quoniam ad ea quae

istered unto my necessities, and to them that were with me.

mihi opus erant et his qui mecum sunt ministrauerunt ma-

20:35 I have shewed you all things, how that so labouring nus istae 20:35 omnia ostendi uobis quoniam sic laboran-

ye ought to support the weak, and to remember the words of the Lord Jesus, how he said, It is more blessed to give

than to receive. tes oportet suscipere infirmos ac meminisse uerbi Domini

Iesu quoniam ipse dixit beatius est magis dare quam acci-

20:36 And when he had thus spoken, he kneeled down, pere 20:36 et cum haec dixisset positis genibus suis cum

and prayed with them all.

20:37 And they all wept sore, and fell on Paul’s neck, and omnibus illis orauit 20:37 magnus autem fletus factus est

kissed him,

omnium et procumbentes super collum Pauli osculabantur

20:38 Sorrowing most of all for the words which he spake, eum 20:38 dolentes maxime in uerbo quo dixerat quoniam that they should see his face no more. And they accompanied him unto the ship.

Actus Apostolorum

amplius faciem eius non essent uisuri et deducebant eum ad nauem

21:1 cum autem factum esset ut nauigaremus abstracti ab 21:1 And it came to pass, that after we were gotten from

them, and had launched, we came with a straight course unto Coos, and the day following unto Rhodes, and from

eis recto cursu uenimus Cho et sequenti die Rhodum et in- thence unto Patara: de Patara 21:2 et cum inuenissemus nauem transfretantem 21:2 And finding a ship sailing over unto Phenicia, we

went aboard, and set forth.

in Foenicen ascendentes nauigauimus 21:3 cum paruisse- 21:3 Now when we had discovered Cyprus, we left it on

the left hand, and sailed into Syria, and landed at Tyre: for there the ship was to unlade her burden.

mus autem Cypro et relinquentes eam ad sinistram nauigabamus in Syriam et uenimus Tyrum ibi enim nauis erat expositura onus 21:4 inuentis autem discipulis mansimus ibi 21:4 And finding disciples, we tarried there seven days:

who said to Paul through the Spirit, that he should not go up to Jerusalem.

diebus septem qui Paulo dicebant per Spiritum ne ascenderet Hierosolymam 21:5 et explicitis diebus profecti ibamus 21:5 And when we had accomplished those days, we de-

parted and went our way; and they all brought us on our way, with wives and children, till we were out of the city:

deducentibus nos omnibus cum uxoribus et filiis usque fo- and we kneeled down on the shore, and prayed. ras ciuitatem et positis genibus in litore orauimus 21:6 et 21:6 And when we had taken our leave one of another, we

took ship; and they returned home again.

cum ualefecissemus inuicem ascendimus in nauem illi autem rediirunt in sua 21:7 nos uero nauigatione explicita a 21:7 And when we had finished our course from Tyre, we

came to Ptolemais, and saluted the brethren, and abode with them one day.

Tyro descendimus Ptolomaida et salutatis fratribus mansimus die una apud illos 21:8 alia autem die profecti uenimus 21:8 And the next day we that were of Paul’s company

departed, and came unto Caesarea: and we entered into the house of Philip the evangelist, which was one of the

Caesaream et intrantes in domum Philippi euangelistae qui seven; and abode with him.

erat de septem mansimus apud eum 21:9 huic autem erant 21:9 And the same man had four daughters, virgins, which

did prophesy.

filiae quattuor uirgines prophetantes 21:10 et cum morare- 21:10 And as we tarried there many days, there came down

from Judaea a certain prophet, named Agabus.

mur per dies aliquot superuenit quidam a Iudaea propheta nomine Agabus 21:11 is cum uenisset ad nos tulit zonam 21:11 And when he was come unto us, he took Paul’s gir-

dle, and bound his own hands and feet, and said, Thus saith the Holy Ghost, So shall the Jews at Jerusalem bind the man that owneth this girdle, and shall deliver him into the hands of the Gentiles.

Pauli et alligans sibi pedes et manus dixit haec dicit Spiritus Sanctus uirum cuius est zona haec sic alligabunt in Hierusalem Iudaei et tradent in manus gentium 21:12 quod cum au- 21:12 And when we heard these things, both we, and they

of that place, besought him not to go up to Jerusalem.

dissemus rogabamus nos et qui loci illius erant ne ascenderet Hierosolymam 21:13 tunc respondit Paulus et dixit quid fa- 21:13 Then Paul answered, What mean ye to weep and to

break mine heart? for I am ready not to be bound only, but also to die at Jerusalem for the name of the Lord Jesus.

citis flentes et adfligentes cor meum ego enim non solum alligari sed et mori in Hierusalem paratus sum propter nomen Domini Iesu 21:14 et cum ei suadere non possemus 21:14 And when he would not be persuaded, we ceased,

saying, The will of the Lord be done.

quieuimus dicentes Domini uoluntas fiat 21:15 post dies au- 21:15 And after those days we took up our carriages, and

went up to Jerusalem.

tem istos praeparati ascendebamus Hierusalem 21:16 uene- 21:16 There went with us also certain of the disciples of

Caesarea, and brought with them one Mnason of Cyprus, an old disciple, with whom we should lodge.

runt autem et ex discipulis a Caesarea nobiscum adducen-

Biblia Sacra Vulgata

tes apud quem hospitaremur Mnasonem quendam Cyprium

21:17 And when we were come to Jerusalem, the brethren antiquum discipulum 21:17 et cum uenissemus Hierosoly-

received us gladly.

21:18 And the day following Paul went in with us unto mam libenter exceperunt nos fratres 21:18 sequenti autem

James; and all the elders were present.

die introibat Paulus nobiscum ad Iacobum omnesque col-

21:19 And when he had saluted them, he declared particu- lecti sunt seniores 21:19 quos cum salutasset narrabat per

larly what things God had wrought among the Gentiles by his ministry.

singula quae fecisset Deus in gentibus per ministerium ip-

21:20 And when they heard it, they glorified the Lord, and sius 21:20 at illi cum audissent magnificabant Deum dixe-

said unto him, Thou seest, brother, how many thousands of Jews there are which believe; and they are all zealous

of the law: runtque ei uides frater quot milia sint in Iudaeis qui credi-21:21 And they are informed of thee, that thou teachest all derunt et omnes aemulatores sunt legis 21:21 audierunt au-

the Jews which are among the Gentiles to forsake Moses, saying that they ought not to circumcise their children, nei-

ther to walk after the customs. tem de te quia discessionem doceas a Mose eorum qui per

gentes sunt Iudaeorum dicens non debere circumcidere eos

21:22 What is it therefore? the multitude must needs come filios suos neque secundum consuetudinem ingredi 21:22 together: for they will hear that thou art come.

quid ergo est utique oportet conuenire multitudinem audient

21:23 Do therefore this that we say to thee: We have four enim te superuenisse 21:23 hoc ergo fac quod tibi dicimus

men which have a vow on them;

21:24 Them take, and purify thyself with them, and be at sunt nobis uiri quattuor uotum habentes super se 21:24 his

charges with them, that they may shave their heads: and all may know that those things, whereof they were informed concerning thee, are nothing; but that thou thyself also walkest orderly, and keepest the law.

adsumptis sanctifica te cum illis et inpende in illis ut radant capita et scient omnes quia quae de te audierunt falsa sunt

21:25 As touching the Gentiles which believe, we have sed ambulas et ipse custodiens legem 21:25 de his autem

written and concluded that they observe no such thing, save only that they keep themselves from things offered to idols, and from blood, and from strangled, and from fornication.

qui crediderunt ex gentibus nos scripsimus iudicantes ut abstineant se ab idolis immolato et sanguine et suffocato et

21:26 Then Paul took the men, and the next day purifying fornicatione 21:26 tunc Paulus adsumptis uiris postera die

himself with them entered into the temple, to signify the accomplishment of the days of purification, until that an

offering should be offered for every one of them. purificatus cum illis intrauit in templum adnuntians exple-

tionem dierum purificationis donec offerretur pro unoquo-

21:27 And when the seven days were almost ended, the que eorum oblatio 21:27 dum autem septem dies consum-

Jews which were of Asia, when they saw him in the temple, stirred up all the people, and laid hands on him,

marentur hii qui de Asia erant Iudaei cum uidissent eum in templo concitauerunt omnem populum et iniecerunt ei ma-

21:28 Crying out, Men of Israel, help: This is the man, nus clamantes 21:28 uiri israhelitae adiuuate hic est homo

that teacheth all men every where against the people, and the law, and this place: and further brought Greeks also

into the temple, and hath polluted this holy place. qui aduersus populum et legem et locum hunc omnes ubi-

que docens insuper et gentiles induxit in templum et uiola-

21:29 (For they had seen before with him in the city uit sanctum locum istum 21:29 uiderant enim Trophimum

Trophimus an Ephesian, whom they supposed that Paul had brought into the temple.)

Ephesium in ciuitate cum ipso quem aestimauerunt quoniam

21:30 And all the city was moved, and the people ran to- in templum induxisset Paulus 21:30 commotaque est ciuitas

gether: and they took Paul, and drew him out of the temple: and forthwith the doors were shut.

tota et facta est concursio populi et adprehendentes Paulum

Actus Apostolorum

trahebant eum extra templum et statim clausae sunt ianuae 21:31 quaerentibus autem eum occidere nuntiatum est tribuno 21:31 And as they went about to kill him, tidings came

unto the chief captain of the band, that all Jerusalem was in an uproar.

cohortis quia tota confunditur Hierusalem 21:32 qui statim

21:32 Who immediately took soldiers and centurions, and ran down unto them: and when they saw the chief captain and the soldiers, they left beating of Paul.

adsumptis militibus et centurionibus decucurrit ad illos qui cum uidissent tribunum et milites cessauerunt percutere Paulum 21:33 tunc accedens tribunus adprehendit eum et iussit 21:33 Then the chief captain came near, and took him,

and commanded him to be bound with two chains; and demanded who he was, and what he had done.

alligari catenis duabus et interrogabat quis esset et quid fecisset 21:34 alii autem aliud clamabant in turba et cum non 21:34 And some cried one thing, some another, among

the multitude: and when he could not know the certainty for the tumult, he commanded him to be carried into the

posset certum cognoscere prae tumultu iussit duci eum in castle. castra 21:35 et cum uenisset ad gradus contigit ut portare- 21:35 And when he came upon the stairs, so it was, that he

was borne of the soldiers for the violence of the people.

tur a militibus propter uim populi 21:36 sequebatur enim 21:36 For the multitude of the people followed after, cry-

ing, Away with him.

multitudo populi clamans tolle eum 21:37 et cum coepis- 21:37 And as Paul was to be led into the castle, he said

unto the chief captain, May I speak unto thee? Who said, Canst thou speak Greek?

set induci in castra Paulus dicit tribuno si licet mihi loqui

aliquid ad te qui dixit graece nosti 21:38 nonne tu es Ae- 21:38 Art not thou that Egyptian, which before these days madest an uproar, and leddest out into the wilderness four thousand men that were murderers?

gyptius qui ante hos dies tumultum concitasti et eduxisti in desertum quattuor milia uirorum sicariorum 21:39 et dixit 21:39 But Paul said, I am a man which am a Jew of Tarsus,

a city in Cilicia, a citizen of no mean city: and, I beseech thee, suffer me to speak unto the people.

ad eum Paulus ego homo sum quidem iudaeus a Tarso Ciliciae non ignotae ciuitatis municeps rogo autem te permitte mihi loqui ad populum 21:40 et cum ille permisisset Paulus 21:40 And when he had given him licence, Paul stood on

the stairs, and beckoned with the hand unto the people. And when there was made a great silence, he spake unto

stans in gradibus annuit manu ad plebem et magno silentio them in the Hebrew tongue, saying, facto adlocutus est hebraea lingua dicens

22:1 uiri fratres et patres audite quam ad uos nunc reddo 22:1 Men, brethren, and fathers, hear ye my defence which

I make now unto you.

rationem 22:2 cum audissent autem quia hebraea lingua lo- 22:2 (And when they heard that he spake in the Hebrew

tongue to them, they kept the more silence: and he saith,)

quitur ad illos magis praestiterunt silentium 22:3 et dixit 22:3 I am verily a man which am a Jew, born in Tarsus,

a city in Cilicia, yet brought up in this city at the feet of Gamaliel, and taught according to the perfect manner of the law of the fathers, and was zealous toward God, as ye all are this day.

ego sum uir iudaeus natus Tarso Ciliciae nutritus autem in ista ciuitate secus pedes Gamalihel eruditus iuxta ueritatem paternae legis aemulator legis sicut et uos omnes estis ho-

die 22:4 qui hanc uiam persecutus sum usque ad mortem 22:4 And I persecuted this way unto the death, binding and

delivering into prisons both men and women.

alligans et tradens in custodias uiros ac mulieres 22:5 sic- 22:5 As also the high priest doth bear me witness, and

all the estate of the elders: from whom also I received letters unto the brethren, and went to Damascus, to bring them which were there bound unto Jerusalem, for to be punished.

ut princeps sacerdotum testimonium mihi reddit et omnes maiores natu a quibus et epistulas accipiens ad fratres Damascum pergebam ut adducerem inde uinctos in Hierusalem uti punirentur 22:6 factum est autem eunte me et adpropin- 22:6 And it came to pass, that, as I made my journey, and

was come nigh unto Damascus about noon, suddenly there shone from heaven a great light round about me.

Biblia Sacra Vulgata

quante Damasco media die subito de caelo circumfulsit me

22:7 And I fell unto the ground, and heard a voice saying lux copiosa 22:7 et decidens in terram audiui uocem dicen-

unto me, Saul, Saul, why persecutest thou me?

22:8 And I answered, Who art thou, Lord? And he said tem mihi Saule Saule quid me persequeris 22:8 ego autem

unto me, I am Jesus of Nazareth, whom thou persecutest.

respondi quis es Domine dixitque ad me ego sum Iesus Na-

22:9 And they that were with me saw indeed the light, and zarenus quem tu persequeris 22:9 et qui mecum erant luwere afraid; but they heard not the voice of him that spake to me.

men quidem uiderunt uocem autem non audierunt eius qui

22:10 And I said, What shall I do, LORD? And the Lord loquebatur mecum 22:10 et dixi quid faciam Domine Do-

said unto me, Arise, and go into Damascus; and there it shall be told thee of all things which are appointed for thee

to do. minus autem dixit ad me surgens uade Damascum et ibi tibi

22:11 And when I could not see for the glory of that light, dicetur de omnibus quae te oporteat facere 22:11 et cum being led by the hand of them that were with me, I came into Damascus.

non uiderem prae claritate luminis illius ad manum deduc-

22:12 And one Ananias, a devout man according to the tus a comitibus ueni Damascum 22:12 Ananias autem qui-

law, having a good report of all the Jews which dwelt there,

dam uir secundum legem testimonium habens ab omnibus

22:13 Came unto me, and stood, and said unto me, Brother habitantibus Iudaeis 22:13 ueniens ad me et adstans dixit

Saul, receive thy sight. And the same hour I looked up upon him.

mihi Saule frater respice et ego eadem hora respexi in eum

22:14 And he said, The God of our fathers hath chosen 22:14 at ille dixit Deus patrum nostrorum praeordinauit te ut

thee, that thou shouldest know his will, and see that Just One, and shouldest hear the voice of his mouth.

cognosceres uoluntatem eius et uideres Iustum et audires uo-

22:15 For thou shalt be his witness unto all men of what cem ex ore eius 22:15 quia eris testis illius ad omnes homi-

thou hast seen and heard.

22:16 And now why tarriest thou? arise, and be baptized, nes eorum quae uidisti et audisti 22:16 et nunc quid moraris

and wash away thy sins, calling on the name of the Lord.

exsurge baptizare et ablue peccata tua inuocato nomine ip-

22:17 And it came to pass, that, when I was come again sius 22:17 factum est autem reuertenti mihi in Hierusalem

to Jerusalem, even while I prayed in the temple, I was in a trance;

et oranti in templo fieri me in stupore mentis 22:18 et ui-

22:18 And saw him saying unto me, Make haste, and get thee quickly out of Jerusalem: for they will not receive thy testimony concerning me.

dere illum dicentem mihi festina et exi uelociter ex Hierusa-

22:19 And I said, Lord, they know that I imprisoned and lem quoniam non recipient testimonium tuum de me 22:19

beat in every synagogue them that believed on thee:

et ego dixi Domine ipsi sciunt quia ego eram concludens in carcerem et caedens per synagogas eos qui credebant in

22:20 And when the blood of thy martyr Stephen was te 22:20 et cum funderetur sanguis Stephani testis tui ego

shed, I also was standing by, and consenting unto his death, and kept the raiment of them that slew him.

adstabam et consentiebam et custodiebam uestimenta inter-

22:21 And he said unto me, Depart: for I will send thee ficientium illum 22:21 et dixit ad me uade quoniam ego in

far hence unto the Gentiles.

22:22 And they gave him audience unto this word, and nationes longe mittam te 22:22 audiebant autem eum usque

then lifted up their voices, and said, Away with such a fellow from the earth: for it is not fit that he should live.

ad hoc uerbum et leuauerunt uocem suam dicentes tolle de

22:23 And as they cried out, and cast off their clothes, and terra eiusmodi non enim fas est eum uiuere 22:23 uocifethrew dust into the air,

rantibus autem eis et proicientibus uestimenta sua et pulue-

22:24 The chief captain commanded him to be brought rem iactantibus in aerem 22:24 iussit tribunus induci eum in

into the castle, and bade that he should be examined by scourging; that he might know wherefore they cried so against him.

Actus Apostolorum

castra et flagellis caedi et torqueri eum ut sciret propter quam causam sic adclamarent ei 22:25 et cum adstrinxissent eum 22:25 And as they bound him with thongs, Paul said unto

the centurion that stood by, Is it lawful for you to scourge a man that is a Roman, and uncondemned?

loris dixit adstanti sibi centurioni Paulus si hominem romanum et indemnatum licet uobis flagellare 22:26 quo audito 22:26 When the centurion heard that, he went and told the

chief captain, saying, Take heed what thou doest: for this man is a Roman.

centurio accessit ad tribunum et nuntiauit dicens quid ac-

turus es hic enim homo ciuis romanus est 22:27 accedens 22:27 Then the chief captain came, and said unto him, Tell

me, art thou a Roman? He said, Yea.

autem tribunus dixit illi dic mihi tu Romanus es at ille dixit etiam 22:28 et respondit tribunus ego multa summa ciui- 22:28 And the chief captain answered, With a great sum

obtained I this freedom. And Paul said, But I was free born.

tatem hanc consecutus sum et Paulus ait ego autem et natus sum 22:29 protinus ergo discesserunt ab illo qui eum torturi 22:29 Then straightway they departed from him which

should have examined him: and the chief captain also was afraid, after he knew that he was a Roman, and because he

erant tribunus quoque timuit postquam resciuit quia ciuis ro- had bound him. manus esset et quia alligasset eum 22:30 postera autem die 22:30 On the morrow, because he would have known the

certainty wherefore he was accused of the Jews, he loosed him from his bands, and commanded the chief priests and all their council to appear, and brought Paul down, and set him before them.

uolens scire diligentius qua ex causa accusaretur a Iudaeis soluit eum et iussit sacerdotes conuenire et omne concilium et producens Paulum statuit inter illos

23:1 intendens autem concilium Paulus ait uiri fratres ego 23:1 And Paul, earnestly beholding the council, said, Men

and brethren, I have lived in all good conscience before God until this day.

omni conscientia bona conuersatus sum ante Deum usque in hodiernum diem 23:2 princeps autem sacerdotum Ana- 23:2 And the high priest Ananias commanded them that

stood by him to smite him on the mouth.

nias praecepit adstantibus sibi percutere os eius 23:3 tunc 23:3 Then said Paul unto him, God shall smite thee, thou

whited wall: for sittest thou to judge me after the law, and commandest me to be smitten contrary to the law?

Paulus ad eum dixit percutiet te Deus paries dealbate et tu sedens iudicas me secundum legem et contra legem iubes

me percuti 23:4 et qui adstabant dixerunt summum sacer- 23:4 And they that stood by said, Revilest thou God’s high

priest?

dotem Dei maledicis 23:5 dixit autem Paulus nesciebam 23:5 Then said Paul, I wist not, brethren, that he was the high priest: for it is written, Thou shalt not speak evil of the ruler of thy people.

fratres quia princeps est sacerdotum scriptum est enim principem populi tui non maledices 23:6 sciens autem Paulus 23:6 But when Paul perceived that the one part were Sad-

ducees, and the other Pharisees, he cried out in the council, Men and brethren, I am a Pharisee, the son of a Pharisee: of the hope and resurrection of the dead I am called in question.

quia una pars esset Sadducaeorum et altera Pharisaeorum exclamauit in concilio uiri fratres ego Pharisaeus sum filius Pharisaeorum de spe et resurrectione mortuorum ego iudicor 23:7 et cum haec dixisset facta est dissensio inter 23:7 And when he had so said, there arose a dissension be-

tween the Pharisees and the Sadducees: and the multitude was divided.

Pharisaeos et Sadducaeos et soluta est multitudo 23:8 Sad-

23:8 For the Sadducees say that there is no resurrection, neither angel, nor spirit: but the Pharisees confess both.

ducaei enim dicunt non esse resurrectionem neque angelum neque spiritum Pharisaei autem utrumque confitentur 23:9 23:9 And there arose a great cry: and the scribes that were

of the Pharisees’ part arose, and strove, saying, We find no evil in this man: but if a spirit or an angel hath spoken to

factus est autem clamor magnus et surgentes quidam Pha- him, let us not fight against God.

Biblia Sacra Vulgata

risaeorum pugnabant dicentes nihil mali inuenimus in ho-

23:10 And when there arose a great dissension, the chief mine isto quod si spiritus locutus est ei aut angelus 23:10

captain, fearing lest Paul should have been pulled in pieces of them, commanded the soldiers to go down, and to take him by force from among them, and to bring him into the castle.

et cum magna dissensio facta esset timens tribunus ne discerperetur Paulus ab ipsis iussit milites descendere et rapere

23:11 And the night following the Lord stood by him, and eum de medio eorum ac deducere eum in castra 23:11 sesaid, Be of good cheer, Paul: for as thou hast testified of me in Jerusalem, so must thou bear witness also at Rome.

quenti autem nocte adsistens ei Dominus ait constans esto sicut enim testificatus es de me Hierusalem sic te oportet et

23:12 And when it was day, certain of the Jews banded Romae testificari 23:12 facta autem die collegerunt se qui-

together, and bound themselves under a curse, saying that they would neither eat nor drink till they had killed Paul.

dam ex Iudaeis et deuouerunt se dicentes neque manduca-

23:13 And they were more than forty which had made this turos neque bibituros donec occiderent Paulum 23:13 erant

conspiracy.

autem plus quam quadraginta qui hanc coniurationem fece-

23:14 And they came to the chief priests and elders, and rant 23:14 qui accesserunt ad principes sacerdotum et se-

said, We have bound ourselves under a great curse, that we will eat nothing until we have slain Paul.

niores et dixerunt deuotione deuouimus nos nihil gustaturos

23:15 Now therefore ye with the council signify to the donec occidamus Paulum 23:15 nunc ergo uos notum fachief captain that he bring him down unto you to morrow, as though ye would enquire something more perfectly concerning him: and we, or ever he come near, are ready to kill him.

cite tribuno cum concilio ut producat illum ad uos tamquam aliquid certius cognituri de eo nos uero priusquam adpro-

23:16 And when Paul’s sister’s son heard of their lying in piet parati sumus interficere illum 23:16 quod cum audisset

wait, he went and entered into the castle, and told Paul.

filius sororis Pauli insidias uenit et intrauit in castra nun-

23:17 Then Paul called one of the centurions unto him, tiauitque Paulo 23:17 uocans autem Paulus ad se unum ex

and said, Bring this young man unto the chief captain: for he hath a certain thing to tell him.

centurionibus ait adulescentem hunc perduc ad tribunum ha-

23:18 So he took him, and brought him to the chief cap- bet enim aliquid indicare illi 23:18 et ille quidem adsumens

tain, and said, Paul the prisoner called me unto him, and prayed me to bring this young man unto thee, who hath

something to say unto thee. eum duxit ad tribunum et ait uinctus Paulus uocans rogauit

me hunc adulescentem perducere ad te habentem aliquid lo-

23:19 Then the chief captain took him by the hand, and qui tibi 23:19 adprehendens autem tribunus manum illius

went with him aside privately, and asked him, What is that thou hast to tell me?

secessit cum eo seorsum et interrogauit illum quid est quod

23:20 And he said, The Jews have agreed to desire thee habes indicare mihi 23:20 ille autem dixit Iudaeis conuenit

that thou wouldest bring down Paul to morrow into the council, as though they would enquire somewhat of him

more perfectly. rogare te ut crastina die Paulum producas in concilium quasi 23:21 But do not thou yield unto them: for there lie in aliquid certius inquisituri sint de illo 23:21 tu uero ne credi-

wait for him of them more than forty men, which have bound themselves with an oath, that they will neither eat nor drink till they have killed him: and now are they ready, looking for a promise from thee.

deris illis insidiantur enim ei ex eis uiri amplius quadraginta qui se deuouerunt non manducare neque bibere donec interficiant eum et nunc parati sunt expectantes promissum tuum

23:22 So the chief captain then let the young man de- 23:22 tribunus igitur dimisit adulescentem praecipiens ne cui

part, and charged him, See thou tell no man that thou hast shewed these things to me.

loqueretur quoniam haec nota sibi fecisset 23:23 et uoca-

23:23 And he called unto him two centurions, saying, Make ready two hundred soldiers to go to Caesarea, and horsemen threescore and ten, and spearmen two hundred, at the third hour of the night;

Actus Apostolorum

tis duobus centurionibus dixit illis parate milites ducentos ut eant usque Caesaream et equites septuaginta et lancearios ducentos a tertia hora noctis 23:24 et iumenta praeparate 23:24 And provide them beasts, that they may set Paul on,

and bring him safe unto Felix the governor.

ut inponentes Paulum saluum perducerent ad Felicem praesidem 23:25 23:26 scribens epistulam continentem haec 23:25 And he wrote a letter after this manner:

23:26 Claudius Lysias unto the most excellent governor Felix sendeth greeting.

Claudius Lysias optimo praesidi Felici salutem 23:27 ui-

23:27 This man was taken of the Jews, and should have been killed of them: then came I with an army, and rescued

rum hunc conprehensum a Iudaeis et incipientem interfici ab ^{him, having understood that he was a Roman.} eis superueniens cum exercitu eripui cognito quia Romanus est 23:28 uolensque scire causam quam obiciebant illi de- 23:28 And when I would have known the cause wherefore

they accused him, I brought him forth into their council:

duxi eum in concilium eorum 23:29 quem inueni accusari 23:29 Whom I perceived to be accused of questions of

their law, but to have nothing laid to his charge worthy of death or of bonds.

de quaestionibus legis ipsorum nihil uero dignum morte aut uinculis habentem crimen 23:30 et cum mihi perlatum es- 23:30 And when it was told me how that the Jews laid

wait for the man, I sent straightway to thee, and gave commandment to his accusers also to say before thee what they

set de insidiis quas parauerunt ei misi ad te denuntians et ac- had against him. Farewell. cusatoribus ut dicant apud te 23:31 milites ergo secundum 23:31 Then the soldiers, as it was commanded them, took

Paul, and brought him by night to Antipatris.

praeceptum sibi adsumentes Paulum duxerunt per noctem in Antipatridem 23:32 et postera die dimissis equitibus ut 23:32 On the morrow they left the horsemen to go with

him, and returned to the castle:

irent cum eo reuersi sunt ad castra 23:33 qui cum uenissent 23:33 Who, when they came to Caesarea and delivered the

epistle to the governor, presented Paul also before him.

Caesaream et tradidissent epistulam praesidi statuerunt ante illum et Paulum 23:34 cum legisset autem et interrogasset 23:34 And when the governor had read the letter, he asked

of what province he was. And when he understood that he was of Cilicia;

de qua prouincia esset et cognoscens quia de Cilicia 23:35

23:35 I will hear thee, said he, when thine accusers are also come. And he commanded him to be kept in Herod’s judgment hall.

audiam te inquit cum et accusatores tui uenerint iussitque in praetorio Herodis custodiri eum

24:1 post quinque autem dies descendit princeps sacerdo- 24:1 And after five days Ananias the high priest descended

with the elders, and with a certain orator named Tertullus, who informed the governor against Paul.

tum Ananias cum senioribus quibusdam et Tertullo quodam oratore qui adierunt praesidem aduersus Paulum 24:2 et ci- 24:2 And when he was called forth, Tertullus began to ac-

cuse him, saying, Seeing that by thee we enjoy great quietness, and that very worthy deeds are done unto this nation

tato Paulo coepit accusare Tertullus dicens cum in multa by thy providence, pace agamus per te et multa corrigantur per tuam prouidentiam 24:3 semper et ubique suscipimus optime Felix cum 24:3 We accept it always, and in all places, most noble

Felix, with all thankfulness.

omni gratiarum actione 24:4 ne diutius autem te protraham 24:4 Notwithstanding, that I be not further tedious unto

thee, I pray thee that thou wouldest hear us of thy clemency a few words.

oro breuiter audias nos pro tua clementia 24:5 inuenimus

24:5 For we have found this man a pestilent fellow, and a mover of sedition among all the Jews throughout the world, and a ringleader of the sect of the Nazarenes:

hunc hominem pestiferum et concitantem seditiones omnibus Iudaeis in uniuerso orbe et auctorem seditionis sectae Nazarenorum 24:6 qui etiam templum uiolare conatus est 24:6 Who also hath gone about to profane the temple:

whom we took, and would have judged according to our law.

Biblia Sacra Vulgata

24:7 But the chief captain Lysias came upon us, and with quem et adprehendimus 24:7 24:8 a quo poteris ipse iugreat violence took him away out of our hands, 24:8 Commanding his accusers to come unto thee: by examining of whom thyself mayest take knowledge of all these things, whereof we accuse him.

dicans de omnibus istis cognoscere de quibus nos accusa-

24:9 And the Jews also assented, saying that these things mus eum 24:9 adiecerunt autem et Iudaei dicentes haec ita were so.

24:10 Then Paul, after that the governor had beckoned se habere 24:10 respondit autem Paulus annuente sibi prae-

unto him to speak, answered, Forasmuch as I know that thou hast been of many years a judge unto this nation, I do

the more cheerfully answer for myself: side dicere ex multis annis esse te iudicem genti huic sciens

24:11 Because that thou mayest understand, that there are bono animo pro me satisfaciam 24:11 potes enim cognosyet but twelve days since I went up to Jerusalem for to worship.

cere quia non plus sunt dies mihi quam duodecim ex quo

24:12 And they neither found me in the temple disputing ascendi adorare in Hierusalem 24:12 et neque in templo in-

with any man, neither raising up the people, neither in the synagogues, nor in the city:

uenerunt me cum aliquo disputantem aut concursum facien-

24:13 Neither can they prove the things whereof they now tem turbae neque in synagogis neque in ciuitate 24:13 ne-

accuse me. 24:14 But this I confess unto thee, that after the way which que probare possunt tibi de quibus nunc accusant me 24:14 they call heresy, so worship I the God of my fathers, believing all things which are written in the law and in the

prophets: confiteor autem hoc tibi quod secundum sectam quam di-

cunt heresim sic deseruio patrio Deo meo credens omnibus

24:15 And have hope toward God, which they themselves quae in lege et prophetis scripta sunt 24:15 spem habens in

also allow, that there shall be a resurrection of the dead, both of the just and unjust.

Deum quam et hii ipsi expectant resurrectionem futuram iu-

24:16 And herein do I exercise myself, to have always a storum et iniquorum 24:16 in hoc et ipse studeo sine offen-

conscience void to offence toward God, and toward men.

diculo conscientiam habere ad Deum et ad homines semper

24:17 Now after many years I came to bring alms to my 24:17 post annos autem plures elemosynas facturus in gen-

nation, and offerings.

24:18 Whereupon certain Jews from Asia found me puri- tem meam ueni et oblationes et uota 24:18 in quibus inue-

fied in the temple, neither with multitude, nor with tumult.

nerunt me purificatum in templo non cum turba neque cum

24:19 Who ought to have been here before thee, and ob- tumultu 24:19 quidam autem ex Asia Iudaei quos oportebat

ject, if they had ought against me.

apud te praesto esse et accusare si quid haberent aduersum

24:20 Or else let these same here say, if they have found me 24:20 aut hii ipsi dicant si quid inuenerunt in me iniqui-

any evil doing in me, while I stood before the council, 24:21 Except it be for this one voice, that I cried standing tatis cum stem in concilio 24:21 nisi de una hac solummodo among them, Touching the resurrection of the dead I am called in question by you this day.

uoce qua clamaui inter eos stans quoniam de resurrectione

24:22 And when Felix heard these things, having more mortuorum ego iudicor hodie a uobis 24:22 distulit autem

perfect knowledge of that way, he deferred them, and said, When Lysias the chief captain shall come down, I will

know the uttermost of your matter. illos Felix certissime sciens de uia dicens cum tribunus Ly-24:23 And he commanded a centurion to keep Paul, and sias descenderit audiam uos 24:23 iussitque centurioni cu-

to let him have liberty, and that he should forbid none of his acquaintance to minister or come unto him.

stodiri eum et habere requiem nec quemquam prohibere de

24:24 And after certain days, when Felix came with his suis ministrare ei 24:24 post aliquot autem dies ueniens Fe-

wife Drusilla, which was a Jewess, he sent for Paul, and heard him concerning the faith in Christ.

lix cum Drusilla uxore sua quae erat Iudaea uocauit Paulum

24:25 And as he reasoned of righteousness, temperance, et audiuit ab eo fidem quae est in Iesum Christum 24:25 and judgment to come, Felix trembled, and answered, Go thy way for this time; when I have a convenient season, I

will call for thee. disputante autem illo de iustitia et castitate et de iudicio fu-

Actus Apostolorum

turo timefactus Felix respondit quod nunc adtinet uade tempore autem oportuno accersiam te 24:26 simul et sperans 24:26 He hoped also that money should have been given

him of Paul, that he might loose him: wherefore he sent for him the oftener, and communed with him.

quia pecunia daretur a Paulo propter quod et frequenter ac-

cersiens eum loquebatur cum eo 24:27 biennio autem ex- 24:27 But after two years Porcius Festus came into Felix’ room: and Felix, willing to shew the Jews a pleasure, left Paul bound.

pleto accepit successorem Felix Porcium Festum uolens autem gratiam praestare Iudaeis Felix reliquit Paulum uinctum

25:1 Festus ergo cum uenisset in prouinciam post triduum 25:1 Now when Festus was come into the province, after

three days he ascended from Caesarea to Jerusalem.

ascendit Hierosolymam a Caesarea 25:2 adieruntque eum 25:2 Then the high priest and the chief of the Jews in-

formed him against Paul, and besought him,

principes sacerdotum et primi Iudaeorum aduersus Paulum et rogabant eum 25:3 postulantes gratiam aduersum eum ut 25:3 And desired favour against him, that he would send

for him to Jerusalem, laying wait in the way to kill him.

iuberet perduci eum Hierusalem insidias tendentes ut eum interficerent in uia 25:4 Festus autem respondit seruari Pau- 25:4 But Festus answered, that Paul should be kept at Cae-

sarea, and that he himself would depart shortly thither.

lum in Caesarea se autem maturius profecturum 25:5 qui 25:5 Let them therefore, said he, which among you are

able, go down with me, and accuse this man, if there be any wickedness in him.

ergo in uobis ait potentes sunt descendentes simul si quod est in uiro crimen accusent eum 25:6 demoratus autem inter 25:6 And when he had tarried among them more than ten

days, he went down unto Caesarea; and the next day sitting on the judgment seat commanded Paul to be brought.

eos dies non amplius quam octo aut decem descendit Caesaream et altera die sedit pro tribunali et iussit Paulum adduci 25:7 qui cum perductus esset circumsteterunt eum qui 25:7 And when he was come, the Jews which came down

from Jerusalem stood round about, and laid many and grievous complaints against Paul, which they could not

ab Hierosolyma descenderant Iudaei multas et graues cau- prove. sas obicientes quas non poterant probare 25:8 Paulo autem 25:8 While he answered for himself, Neither against the

law of the Jews, neither against the temple, nor yet against Caesar, have I offended any thing at all.

rationem reddente quoniam neque in legem Iudaeorum neque in templum neque in Caesarem quicquam peccaui 25:9 25:9 But Festus, willing to do the Jews a pleasure, an-

swered Paul, and said, Wilt thou go up to Jerusalem, and there be judged of these things before me?

Festus autem uolens Iudaeis gratiam praestare respondens Paulo dixit uis Hierosolymam ascendere et ibi de his iudicari apud me 25:10 dixit autem Paulus ad tribunal Caesaris 25:10 Then said Paul, I stand at Caesar’s judgment seat,

where I ought to be judged: to the Jews have I done no wrong, as thou very well knowest.

sto ubi me oportet iudicari Iudaeis non nocui sicut tu melius nosti 25:11 si enim nocui aut dignum morte aliquid 25:11 For if I be an offender, or have committed any thing

worthy of death, I refuse not to die: but if there be none of these things whereof these accuse me, no man may deliver

feci non recuso mori si uero nihil est eorum quae hii ac- me unto them. I appeal unto Caesar. cusant me nemo potest me illis donare Caesarem appello 25:12 tunc Festus cum consilio locutus respondit Caesarem 25:12 Then Festus, when he had conferred with the coun-

cil, answered, Hast thou appealed unto Caesar? unto Caesar shalt thou go.

appellasti ad Caesarem ibis 25:13 et cum dies aliquot trans-

25:13 And after certain days king Agrippa and Bernice came unto Caesarea to salute Festus.

acti essent Agrippa rex et Bernice descenderunt Caesaream ad salutandum Festum 25:14 et cum dies plures ibi demo- 25:14 And when they had been there many days, Festus

declared Paul’s cause unto the king, saying, There is a certain man left in bonds by Felix:

Biblia Sacra Vulgata

rarentur Festus regi indicauit de Paulo dicens uir quidam est

25:15 About whom, when I was at Jerusalem, the chief derelictus a Felice uinctus 25:15 de quo cum essem Hie-

priests and the elders of the Jews informed me, desiring to have judgment against him.

rosolymis adierunt me principes sacerdotum et seniores Iu-

25:16 To whom I answered, It is not the manner of the daeorum postulantes aduersus illum damnationem 25:16 ad

Romans to deliver any man to die, before that he which is accused have the accusers face to face, and have licence to answer for himself concerning the crime laid against him. quos respondi quia non est consuetudo Romanis donare ali-

quem hominem priusquam is qui accusatur praesentes habeat accusatores locumque defendendi accipiat ad abluenda

25:17 Therefore, when they were come hither, without any crimina 25:17 cum ergo huc conuenissent sine ulla diladelay on the morrow I sat on the judgment seat, and commanded the man to be brought forth.

tione sequenti die sedens pro tribunali iussi adduci uirum

25:18 Against whom when the accusers stood up, they 25:18 de quo cum stetissent accusatores nullam causam de-

brought none accusation of such things as I supposed:

25:19 But had certain questions against him of their own ferebant de quibus ego suspicabar malum 25:19 quaestio-

superstition, and of one Jesus, which was dead, whom Paul affirmed to be alive.

nes uero quasdam de sua superstitione habebant aduersus eum et de quodam Iesu defuncto quem adfirmabat Paulus

25:20 And because I doubted of such manner of questions, uiuere 25:20 haesitans autem ego de huiusmodi quaestione

I asked him whether he would go to Jerusalem, and there be judged of these matters.

dicebam si uellet ire Hierosolymam et ibi iudicari de istis

25:21 But when Paul had appealed to be reserved unto the 25:21 Paulo autem appellante ut seruaretur ad Augusti cog-

hearing of Augustus, I commanded him to be kept till I might send him to Caesar.

nitionem iussi seruari eum donec mittam eum ad Caesarem

25:22 Then Agrippa said unto Festus, I would also hear 25:22 Agrippa autem ad Festum uolebam et ipse hominem the man myself. To morrow, said he, thou shalt hear him.

25:23 And on the morrow, when Agrippa was come, and audire cras inquit audies eum 25:23 altera autem die cum

Bernice, with great pomp, and was entered into the place of hearing, with the chief captains, and principal men of

the city, at Festus’ commandment Paul was brought forth. uenisset Agrippa et Bernice cum multa ambitione et intro-

issent in auditorium cum tribunis et uiris principalibus ci-

25:24 And Festus said, King Agrippa, and all men which uitatis et iubente Festo adductus est Paulus 25:24 et dixit

are here present with us, ye see this man, about whom all the multitude of the Jews have dealt with me, both at Jerusalem, and also here, crying that he ought not to live any longer.

Festus Agrippa rex et omnes qui simul adestis nobiscum uiri uidetis hunc de quo omnis multitudo Iudaeorum interpellauit me Hierosolymis petens et hic clamantes non oportere eum

25:25 But when I found that he had committed nothing uiuere amplius 25:25 ego uero conperi nihil dignum eum

worthy of death, and that he himself hath appealed to Augustus, I have determined to send him.

morte admisisse ipso autem hoc appellante Augustum iudi-

25:26 Of whom I have no certain thing to write unto my caui mittere 25:26 de quo quid certum scribam domino non

lord. Wherefore I have brought him forth before you, and specially before thee, O king Agrippa, that, after examina-

tion had, I might have somewhat to write. habeo propter quod produxi eum ad uos et maxime ad te rex 25:27 For it seemeth to me unreasonable to send a pris- Agrippa ut interrogatione facta habeam quid scribam 25:27

oner, and not withal to signify the crimes laid against him.

sine ratione enim mihi uidetur mittere uinctum et causas eius non significare

26:1 Then Agrippa said unto Paul, Thou art permitted to 26:1 Agrippa uero ad Paulum ait permittitur tibi loqui pro

speak for thyself. Then Paul stretched forth the hand, and answered for himself:

Actus Apostolorum

temet ipso tunc Paulus extenta manu coepit rationem reddere 26:2 de omnibus quibus accusor a Iudaeis rex Agrippa 26:2 I think myself happy, king Agrippa, because I shall

answer for myself this day before thee touching all the things whereof I am accused of the Jews:

aestimo me beatum apud te cum sim defensurus me hodie

26:3 maxime te sciente omnia quae apud Iudaeos sunt con- 26:3 Especially because I know thee to be expert in all customs and questions which are among the Jews: wherefore I beseech thee to hear me patiently.

suetudines et quaestiones propter quod obsecro patienter me audias 26:4 et quidem uitam meam a iuuentute quae ab in- 26:4 My manner of life from my youth, which was at the

first among mine own nation at Jerusalem, know all the Jews;

itio fuit in gente mea in Hierosolymis nouerunt omnes Iudaei 26:5 praescientes me ab initio si uelint testimonium perhi- 26:5 Which knew me from the beginning, if they would

testify, that after the most straitest sect of our religion I lived a Pharisee.

bere quoniam secundum certissimam sectam nostrae religionis uixi Pharisaeus 26:6 et nunc in spe quae ad patres 26:6 And now I stand and am judged for the hope of the

promise made of God, unto our fathers:

nostros repromissionis facta est a Deo sto iudicio subiectus 26:7 in quam duodecim tribus nostrae nocte ac die deseruien- 26:7 Unto which promise our twelve tribes, instantly serv-

ing God day and night, hope to come. For which hope’s sake, king Agrippa, I am accused of the Jews.

tes sperant deuenire de qua spe accusor a Iudaeis rex 26:8

26:8 Why should it be thought a thing incredible with you, that God should raise the dead?

quid incredibile iudicatur apud uos si Deus mortuos suscitat 26:9 et ego quidem existimaueram me aduersus nomen Iesu 26:9 I verily thought with myself, that I ought to do many

things contrary to the name of Jesus of Nazareth.

Nazareni debere multa contraria agere 26:10 quod et feci 26:10 Which thing I also did in Jerusalem: and many of

the saints did I shut up in prison, having received authority from the chief priests; and when they were put to death, I

Hierosolymis et multos sanctorum ego in carceribus inclusi gave my voice against them. a principibus sacerdotum potestate accepta et cum occiderentur detuli sententiam 26:11 et per omnes synagogas fre- 26:11 And I punished them oft in every synagogue, and

compelled them to blaspheme; and being exceedingly mad against them, I persecuted them even unto strange cities.

quenter puniens eos conpellebam blasphemare et amplius insaniens in eos persequebar usque in exteras ciuitates 26:12 26:12 Whereupon as I went to Damascus with authority

and commission from the chief priests,

in quibus dum irem Damascum cum potestate et permissu principum sacerdotum 26:13 die media in uia uidi rex de 26:13 At midday, O king, I saw in the way a light from

heaven, above the brightness of the sun, shining round about me and them which journeyed with me.

caelo supra splendorem solis circumfulsisse me lumen et eos qui mecum simul erant 26:14 omnesque nos cum decidis- 26:14 And when we were all fallen to the earth, I heard a

voice speaking unto me, and saying in the Hebrew tongue, Saul, Saul, why persecutest thou me? it is hard for thee to

semus in terram audiui uocem loquentem mihi hebraica lin- kick against the pricks. gua Saule Saule quid me persequeris durum est tibi contra stimulum calcitrare 26:15 ego autem dixi quis es Domine 26:15 And I said, Who art thou, Lord? And he said, I am

Jesus whom thou persecutest.

Dominus autem dixit ego sum Iesus quem tu persequeris 26:16 sed exsurge et sta super pedes tuos ad hoc enim ap- 26:16 But rise, and stand upon thy feet: for I have ap-

peared unto thee for this purpose, to make thee a minister and a witness both of these things which thou hast seen,

parui tibi ut constituam te ministrum et testem eorum quae and of those things in the which I will appear unto thee; uidisti et eorum quibus apparebo tibi 26:17 eripiens te de 26:17 Delivering thee from the people, and from the Gen-

tiles, unto whom now I send thee,

populo et gentibus in quas nunc ego mitto te 26:18 aperire 26:18 To open their eyes, and to turn them from darkness

to light, and from the power of Satan unto God, that they may receive forgiveness of sins, and inheritance among them which are sanctified by faith that is in me.

Biblia Sacra Vulgata

oculos eorum ut conuertantur a tenebris ad lucem et de potestate Satanae ad Deum ut accipiant remissionem peccatorum

26:19 Whereupon, O king Agrippa, I was not disobedient et sortem inter sanctos per fidem quae est in me 26:19 unde

unto the heavenly vision:

26:20 But shewed first unto them of Damascus, and at rex Agrippa non fui incredulus caelestis uisionis 26:20 sed

Jerusalem, and throughout all the coasts of Judaea, and then to the Gentiles, that they should repent and turn to

God, and do works meet for repentance. his qui sunt Damasci primum et Hierosolymis et in omnem regionem Iudaeae et gentibus adnuntiabam ut paenitentiam agerent et conuerterentur ad Deum digna paenitentiae opera

26:21 For these causes the Jews caught me in the temple, facientes 26:21 hac ex causa me Iudaei cum essem in temand went about to kill me.

26:22 Having therefore obtained help of God, I continue plo conprehensum temptabant interficere 26:22 auxilio au-

unto this day, witnessing both to small and great, saying none other things than those which the prophets and

Moses did say should come: tem adiutus Dei usque in hodiernum diem sto testificans mi-

nori atque maiori nihil extra dicens quam ea quae prophetae

26:23 That Christ should suffer, and that he should be the sunt locuti futura esse et Moses 26:23 si passibilis Christus

first that should rise from the dead, and should shew light unto the people, and to the Gentiles.

si primus ex resurrectione mortuorum lumen adnuntiaturus

26:24 And as he thus spake for himself, Festus said with est populo et gentibus 26:24 haec loquente eo et rationem

a loud voice, Paul, thou art beside thyself; much learning doth make thee mad.

reddente Festus magna uoce dixit insanis Paule multae te

26:25 But he said, I am not mad, most noble Festus; but litterae ad insaniam conuertunt 26:25 at Paulus non insanio

speak forth the words of truth and soberness.

inquit optime Feste sed ueritatis et sobrietatis uerba eloquor

26:26 For the king knoweth of these things, before whom 26:26 scit enim de his rex ad quem et constanter loquor latere

also I speak freely: for I am persuaded that none of these things are hidden from him; for this thing was not done in

a corner. enim eum nihil horum arbitror neque enim in angulo quic-26:27 King Agrippa, believest thou the prophets? I know quam horum gestum est 26:27 credis rex Agrippa prophetis

that thou believest.

26:28 Then Agrippa said unto Paul, Almost thou per- scio quia credis 26:28 Agrippa autem ad Paulum in modico

suadest me to be a Christian.

26:29 And Paul said, I would to God, that not only thou, suades me Christianum fieri 26:29 et Paulus opto apud De-

but also all that hear me this day, were both almost, and altogether such as I am, except these bonds.

um et in modico et in magno non tantum te sed et omnes hos qui audiunt hodie fieri tales qualis et ego sum exceptis uin-

26:30 And when he had thus spoken, the king rose up, and culis his 26:30 et exsurrexit rex et praeses et Bernice et qui

the governor, and Bernice, and they that sat with them:

26:31 And when they were gone aside, they talked be- adsidebant eis 26:31 et cum secessissent loquebantur ad in-

tween themselves, saying, This man doeth nothing worthy of death or of bonds.

uicem dicentes quia nihil morte aut uinculorum dignum quid

26:32 Then said Agrippa unto Festus, This man might facit homo iste 26:32 Agrippa autem Festo dixit dimitti po-

have been set at liberty, if he had not appealed unto Caesar.

terat homo hic si non appellasset Caesarem

27:1 And when it was determined that we should sail into 27:1 ut autem iudicatum est eum nauigare in Italiam et Italy, they delivered Paul and certain other prisoners unto one named Julius, a centurion of Augustus’ band.

tradi Paulum cum reliquis custodiis centurioni nomine Iu-

27:2 And entering into a ship of Adramyttium, we lio cohortis Augustae 27:2 ascendentes autem nauem ha-

launched, meaning to sail by the coasts of Asia; one Aristarchus, a Macedonian of Thessalonica, being with us.

drumetinam incipientem nauigare circa Asiae loca sustuli-

Actus Apostolorum

mus perseuerante nobiscum Aristarcho Macedone Thessalonicense 27:3 sequenti autem die deuenimus Sidonem hu- 27:3 And the next day we touched at Sidon. And Julius

courteously entreated Paul, and gave him liberty to go unto his friends to refresh himself.

mane autem tractans Iulius Paulum permisit ad amicos ire et curam sui agere 27:4 et inde cum sustulissemus subnaui- 27:4 And when we had launched from thence, we sailed

under Cyprus, because the winds were contrary.

gauimus Cypro propterea quod essent uenti contrarii 27:5 27:5 And when we had sailed over the sea of Cilicia and

Pamphylia, we came to Myra, a city of Lycia.

et pelagus Ciliciae et Pamphiliae nauigantes uenimus Lystram quae est Lyciae 27:6 et ibi inueniens centurio nauem 27:6 And there the centurion found a ship of Alexandria

sailing into Italy; and he put us therein.

alexandrinam nauigantem in Italiam transposuit nos in eam 27:7 et cum multis diebus tarde nauigaremus et uix deue- 27:7 And when we had sailed slowly many days, and

scarce were come over against Cnidus, the wind not suffering us, we sailed under Crete, over against Salmone;

nissemus contra Cnidum prohibente nos uento adnauigauimus Cretae secundum Salmonem 27:8 et uix iuxta nauigan- 27:8 And, hardly passing it, came unto a place which is

called The fair havens; nigh whereunto was the city of Lasea.

tes uenimus in locum quendam qui uocatur Boni portus cui iuxta erat ciuitas Thalassa 27:9 multo autem tempore per- 27:9 Now when much time was spent, and when sailing

was now dangerous, because the fast was now already past, Paul admonished them,

acto et cum iam non esset tuta nauigatio eo quod et ieiunium iam praeterisset consolabatur Paulus 27:10 dicens eis uiri 27:10 And said unto them, Sirs, I perceive that this voyage

will be with hurt and much damage, not only of the lading and ship, but also of our lives.

uideo quoniam cum iniuria et multo damno non solum oneris et nauis sed etiam animarum nostrarum incipit esse nauigatio 27:11 centurio autem gubernatori et nauclerio magis credebat 27:11 Nevertheless the centurion believed the master and

the owner of the ship, more than those things which were spoken by Paul.

quam his quae a Paulo dicebantur 27:12 et cum aptus portus

27:12 And because the haven was not commodious to winter in, the more part advised to depart thence also, if by any means they might attain to Phenice, and there to winter; which is an haven of Crete, and lieth toward the south west and north west.

non esset ad hiemandum plurimi statuerunt consilium nauigare inde si quo modo possent deuenientes Phoenice hiemare portum Cretae respicientem ad africum et ad chorum 27:13 adspirante autem austro aestimantes propositum se te- 27:13 And when the south wind blew softly, supposing

that they had obtained their purpose, loosing thence, they sailed close by Crete.

nere cum sustulissent de Asson legebant Cretam 27:14 non

27:14 But not long after there arose against it a tempestuous wind, called Euroclydon.

post multum autem misit se contra ipsam uentus typhonicus qui uocatur euroaquilo 27:15 cumque arrepta esset na- 27:15 And when the ship was caught, and could not bear

up into the wind, we let her drive.

uis et non posset conari in uentum data naue flatibus ferebamur 27:16 insulam autem quandam decurrentes quae uoca- 27:16 And running under a certain island which is called

Clauda, we had much work to come by the boat:

tur Caudam potuimus uix obtinere scapham 27:17 qua su- 27:17 Which when they had taken up, they used helps,

undergirding the ship; and, fearing lest they should fall into the quicksands, strake sail, and so were driven.

blata adiutoriis utebantur accingentes nauem timentes ne in Syrtim inciderent submisso uase sic ferebantur 27:18 ualide 27:18 And we being exceedingly tossed with a tempest,

the next day they lightened the ship;

autem nobis tempestate iactatis sequenti die iactum fecerunt 27:19 et tertia die suis manibus armamenta nauis proiecerunt 27:19 And the third day we cast out with our own hands

the tackling of the ship.

Biblia Sacra Vulgata

27:20 And when neither sun nor stars in many days ap- 27:20 neque sole autem neque sideribus apparentibus per

peared, and no small tempest lay on us, all hope that we should be saved was then taken away.

plures dies et tempestate non exigua inminente iam ablata

27:21 But after long abstinence Paul stood forth in the erat spes omnis salutis nostrae 27:21 et cum multa ieiuna-

midst of them, and said, Sirs, ye should have hearkened unto me, and not have loosed from Crete, and to have

gained this harm and loss. tio fuisset tunc stans Paulus in medio eorum dixit oportebat

quidem o uiri audito me non tollere a Creta lucrique facere

27:22 And now I exhort you to be of good cheer: for there iniuriam hanc et iacturam 27:22 et nunc suadeo uobis bono

shall be no loss of any man’s life among you, but of the ship.

animo esse amissio enim nullius animae erit ex uobis prae-

27:23 For there stood by me this night the angel of God, terquam nauis 27:23 adstitit enim mihi hac nocte angelus

whose I am, and whom I serve, 27:24 Saying, Fear not, Paul; thou must be brought before Dei cuius sum ego et cui deseruio 27:24 dicens ne timeas Caesar: and, lo, God hath given thee all them that sail with thee.

Paule Caesari te oportet adsistere et ecce donauit tibi Deus

27:25 Wherefore, sirs, be of good cheer: for I believe God, omnes qui nauigant tecum 27:25 propter quod bono animo that it shall be even as it was told me.

estote uiri credo enim Deo quia sic erit quemadmodum dic-

27:26 Howbeit we must be cast upon a certain island. tum est mihi 27:26 in insulam autem quandam oportet nos 27:27 But when the fourteenth night was come, as we were deuenire 27:27 sed posteaquam quartadecima nox superue-

driven up and down in Adria, about midnight the shipmen deemed that they drew near to some country;

nit nauigantibus nobis in Hadria circa mediam noctem suspi-

27:28 And sounded, and found it twenty fathoms: and cabantur nautae apparere sibi aliquam regionem 27:28 qui

when they had gone a little further, they sounded again, and found it fifteen fathoms.

submittentes inuenerunt passus uiginti et pusillum inde se-

27:29 Then fearing lest we should have fallen upon rocks, parati inuenerunt passus quindecim 27:29 timentes autem

they cast four anchors out of the stern, and wished for the day.

ne in aspera loca incideremus de puppi mittentes anchoras

27:30 And as the shipmen were about to flee out of the quattuor optabant diem fieri 27:30 nautis uero quaerentibus

ship, when they had let down the boat into the sea, under colour as though they would have cast anchors out of the

foreship, fugere de naui cum misissent scapham in mare sub obtentu 27:31 Paul said to the centurion and to the soldiers, Except quasi a prora inciperent anchoras extendere 27:31 dixit Pau-

these abide in the ship, ye cannot be saved.

lus centurioni et militibus nisi hii in naui manserint uos salui

27:32 Then the soldiers cut off the ropes of the boat, and fieri non potestis 27:32 tunc absciderunt milites funes sca-

let her fall off.

27:33 And while the day was coming on, Paul besought phae et passi sunt eam excidere 27:33 et cum lux inciperet

them all to take meat, saying, This day is the fourteenth day that ye have tarried and continued fasting, having

taken nothing. fieri rogabat Paulus omnes sumere cibum dicens quartade-

cima hodie die expectantes ieiuni permanetis nihil accipien-

27:34 Wherefore I pray you to take some meat: for this tes 27:34 propter quod rogo uos accipere cibum pro salute

is for your health: for there shall not an hair fall from the head of any of you.

uestra quia nullius uestrum capillus de capite peribit 27:35

27:35 And when he had thus spoken, he took bread, and gave thanks to God in presence of them all: and when he had broken it, he began to eat.

et cum haec dixisset sumens panem gratias egit Deo in con-

27:36 Then were they all of good cheer, and they also took spectu omnium et cum fregisset coepit manducare 27:36 some meat.

animaequiores autem facti omnes et ipsi adsumpserunt ci-

27:37 And we were in all in the ship two hundred three- bum 27:37 eramus uero uniuersae animae in naui ducentae

score and sixteen souls.

Actus Apostolorum

septuaginta sex 27:38 et satiati cibo adleuiabant nauem iac- 27:38 And when they had eaten enough, they lightened the

ship, and cast out the wheat into the sea.

tantes triticum in mare 27:39 cum autem dies factus esset 27:39 And when it was day, they knew not the land: but

they discovered a certain creek with a shore, into the which they were minded, if it were possible, to thrust in the ship.

terram non agnoscebant sinum uero quendam considerabant habentem litus in quem cogitabant si possent eicere nauem 27:40 et cum anchoras abstulissent committebant se mari si- 27:40 And when they had taken up the anchors, they com-

mitted themselves unto the sea, and loosed the rudder bands, and hoised up the mainsail to the wind, and made

mul laxantes iuncturas gubernaculorum et leuato artemone toward shore. secundum flatum aurae tendebant ad litus 27:41 et cum in- 27:41 And falling into a place where two seas met, they

ran the ship aground; and the forepart stuck fast, and remained unmoveable, but the hinder part was broken with

cidissemus in locum bithalassum inpegerunt nauem et prora the violence of the waves. quidem fixa manebat inmobilis puppis uero soluebatur a ui maris 27:42 militum autem consilium fuit ut custodias oc- 27:42 And the soldiers’ counsel was to kill the prisoners,

lest any of them should swim out, and escape.

ciderent ne quis cum enatasset effugeret 27:43 centurio au- 27:43 But the centurion, willing to save Paul, kept them

from their purpose; and commanded that they which could swim should cast themselves first into the sea, and get to

tem uolens seruare Paulum prohibuit fieri iussitque eos qui land: possent natare mittere se primos et euadere et ad terram exire 27:44 et ceteros alios in tabulis ferebant quosdam super 27:44 And the rest, some on boards, and some on broken

pieces of the ship. And so it came to pass, that they escaped all safe to land.

ea quae de naui essent et sic factum est ut omnes animae euaderent ad terram

28:1 et cum euasissemus tunc cognouimus quia Militene 28:1 And when they were escaped, then they knew that the

island was called Melita.

insula uocatur barbari uero praestabant non modicam humanitatem nobis 28:2 accensa enim pyra reficiebant nos om- 28:2 And the barbarous people shewed us no little kind-

ness: for they kindled a fire, and received us every one, because of the present rain, and because of the cold.

nes propter imbrem qui inminebat et frigus 28:3 cum con-

28:3 And when Paul had gathered a bundle of sticks, and laid them on the fire, there came a viper out of the heat, and fastened on his hand.

gregasset autem Paulus sarmentorum aliquantam multitudinem et inposuisset super ignem uipera a calore cum processisset inuasit manum eius 28:4 ut uero uiderunt barbari pen- 28:4 And when the barbarians saw the venomous beast

hang on his hand, they said among themselves, No doubt this man is a murderer, whom, though he hath escaped the

dentem bestiam de manu eius ad inuicem dicebant utique sea, yet vengeance suffereth not to live. homicida est homo hic qui cum euaserit de mari Vltio non sinit uiuere 28:5 et ille quidem excutiens bestiam in ignem 28:5 And he shook off the beast into the fire, and felt no

harm.

nihil mali passus est 28:6 at illi existimabant eum in tumo- 28:6 Howbeit they looked when he should have swollen,

or fallen down dead suddenly: but after they had looked a great while, and saw no harm come to him, they changed

rem conuertendum et subito casurum et mori diu autem illis their minds, and said that he was a god. sperantibus et uidentibus nihil mali in eo fieri conuertentes se dicebant eum esse deum 28:7 in locis autem illis erant 28:7 In the same quarters were possessions of the chief

man of the island, whose name was Publius; who received us, and lodged us three days courteously.

praedia principis insulae nomine Publii qui nos suscipiens triduo benigne exhibuit 28:8 contigit autem patrem Publii 28:8 And it came to pass, that the father of Publius lay sick

of a fever and of a bloody flux: to whom Paul entered in, and prayed, and laid his hands on him, and healed him.

febribus et dysenteria uexatum iacere ad quem Paulus in-

Biblia Sacra Vulgata

trauit et cum orasset et inposuisset ei manus saluauit eum

28:9 So when this was done, others also, which had dis- 28:9 quo facto et omnes qui in insula habebant infirmitates

eases in the island, came, and were healed:

28:10 Who also honoured us with many honours; and accedebant et curabantur 28:10 qui etiam multis honoribus

when we departed, they laded us with such things as were necessary.

nos honorauerunt et nauigantibus inposuerunt quae neces-

28:11 And after three months we departed in a ship of saria erant 28:11 post menses autem tres nauigauimus in

Alexandria, which had wintered in the isle, whose sign was Castor and Pollux.

naue alexandrina quae in insula hiemauerat cui erat insigne

28:12 And landing at Syracuse, we tarried there three Castorum 28:12 et cum uenissemus Syracusam mansimus days.

28:13 And from thence we fetched a compass, and came ibi triduo 28:13 inde circumlegentes deuenimus Regium et

to Rhegium: and after one day the south wind blew, and we came the next day to Puteoli:

post unum diem flante austro secunda die uenimus Puteo-

28:14 Where we found brethren, and were desired to tarry los 28:14 ubi inuentis fratribus rogati sumus manere apud

with them seven days: and so we went toward Rome. 28:15 And from thence, when the brethren heard of us, eos dies septem et sic uenimus Romam 28:15 et inde cum they came to meet us as far as Appii forum, and The three taverns: whom when Paul saw, he thanked God, and took

courage. audissent fratres occurrerunt nobis usque ad Appii Forum et

Tribus Tabernis quos cum uidisset Paulus gratias agens Deo

28:16 And when we came to Rome, the centurion deliv- accepit fiduciam 28:16 cum uenissemus autem Romam per-

ered the prisoners to the captain of the guard: but Paul was suffered to dwell by himself with a soldier that kept

him. missum est Paulo manere sibimet cum custodiente se mi-

28:17 And it came to pass, that after three days Paul called lite 28:17 post tertium autem diem conuocauit primos Iuthe chief of the Jews together: and when they were come together, he said unto them, Men and brethren, though I have committed nothing against the people, or customs of our fathers, yet was I delivered prisoner from Jerusalem into the hands of the Romans.

daeorum cumque conuenissent dicebat eis ego uiri fratres nihil aduersus plebem faciens aut morem paternum uinctus

28:18 Who, when they had examined me, would have let ab Hierosolymis traditus sum in manus Romanorum 28:18

me go, because there was no cause of death in me.

qui cum interrogationem de me habuissent uoluerunt me di-

28:19 But when the Jews spake against it, I was con- mittere eo quod nulla causa esset mortis in me 28:19 con-

strained to appeal unto Caesar; not that I had ought to accuse my nation of.

tradicentibus autem Iudaeis coactus sum appellare Caesarem

28:20 For this cause therefore have I called for you, to see non quasi gentem meam habens aliquid accusare 28:20 proyou, and to speak with you: because that for the hope of Israel I am bound with this chain.

pter hanc igitur causam rogaui uos uidere et adloqui propter

28:21 And they said unto him, We neither received letters spem enim Israhel catena hac circumdatus sum 28:21 at illi

out of Judaea concerning thee, neither any of the brethren that came shewed or spake any harm of thee.

dixerunt ad eum nos neque litteras accepimus de te a Iudaea neque adueniens aliquis fratrum nuntiauit aut locutus

28:22 But we desire to hear of thee what thou thinkest: est quid de te malum 28:22 rogamus autem a te audire quae

for as concerning this sect, we know that every where it is spoken against.

sentis nam de secta hac notum est nobis quia ubique ei con-

28:23 And when they had appointed him a day, there came tradicitur 28:23 cum constituissent autem illi diem uenerunt

many to him into his lodging; to whom he expounded and testified the kingdom of God, persuading them concerning Jesus, both out of the law of Moses, and out of the prophets, from morning till evening.

ad eum in hospitium plures quibus exponebat testificans regnum Dei suadensque eos de Iesu ex lege Mosi et prophe-

28:24 And some believed the things which were spoken, tis a mane usque ad uesperam 28:24 et quidam credebant and some believed not.

Actus Apostolorum

his quae dicebantur quidam uero non credebant 28:25 cum- 28:25 And when they agreed not among themselves, they

departed, after that Paul had spoken one word, Well spake the Holy Ghost by Esaias the prophet unto our fathers,

que inuicem non essent consentientes discedebant dicente Paulo unum uerbum quia bene Spiritus Sanctus locutus est per Esaiam prophetam ad patres nostros 28:26 dicens uade 28:26 Saying, Go unto this people, and say, Hearing ye

shall hear, and shall not understand; and seeing ye shall see, and not perceive:

ad populum istum et dic aure audietis et non intellegetis et uidentes uidebitis et non perspicietis 28:27 incrassatum est 28:27 For the heart of this people is waxed gross, and their

ears are dull of hearing, and their eyes have they closed; lest they should see with their eyes, and hear with their ears, and understand with their heart, and should be converted, and I should heal them.

enim cor populi huius et auribus grauiter audierunt et oculos suos conpresserunt ne forte uideant oculis et auribus audiant et corde intellegant et conuertantur et sanem illos 28:28 no- 28:28 Be it known therefore unto you, that the salvation of

God is sent unto the Gentiles, and that they will hear it.

tum ergo sit uobis quoniam gentibus missum est hoc salutare Dei ipsi et audient 28:29 28:30 mansit autem biennio toto 28:29 And when he had said these words, the Jews de-

parted, and had great reasoning among themselves.

28:30 And Paul dwelt two whole years in his own hired house, and received all that came in unto him,

in suo conducto et suscipiebat omnes qui ingrediebantur ad

eum 28:31 praedicans regnum Dei et docens quae sunt de 28:31 Preaching the kingdom of God, and teaching those things which concern the Lord Jesus Christ, with all confidence, no man forbidding him.

Domino Iesu Christo cum omni fiducia sine prohibitione

Biblia Sacra Vulgata

MDCVI

Epistola Pauli Ad Romanos

1:1 Paulus seruus Christi Iesu uocatus apostolus segrega- 1:1 Paul, a servant of Jesus Christ, called to be an apostle,

separated unto the gospel of God,

tus in euangelium Dei 1:2 quod ante promiserat per prophe- 1:2 (Which he had promised afore by his prophets in the

holy scriptures,)

tas suos in scripturis sanctis 1:3 de Filio suo qui factus est 1:3 Concerning his Son Jesus Christ our Lord, which was

made of the seed of David according to the flesh;

ex semine Dauid secundum carnem 1:4 qui praedestinatus 1:4 And declared to be the Son of God with power, ac-

cording to the spirit of holiness, by the resurrection from the dead:

est Filius Dei in uirtute secundum Spiritum sanctificationis ex resurrectione mortuorum Iesu Christi Domini nostri 1:5 1:5 By whom we have received grace and apostleship, for

obedience to the faith among all nations, for his name:

per quem accepimus gratiam et apostolatum ad oboediendum fidei in omnibus gentibus pro nomine eius 1:6 in qui- 1:6 Among whom are ye also the called of Jesus Christ: bus estis et uos uocati Iesu Christi 1:7 omnibus qui sunt 1:7 To all that be in Rome, beloved of God, called to be

saints: Grace to you and peace from God our Father, and the Lord Jesus Christ.

Romae dilectis Dei uocatis sanctis gratia uobis et pax a Deo

Patre nostro et Domino Iesu Christo 1:8 primum quidem 1:8 First, I thank my God through Jesus Christ for you all,

that your faith is spoken of throughout the whole world.

gratias ago Deo meo per Iesum Christum pro omnibus uobis quia fides uestra adnuntiatur in uniuerso mundo 1:9 te- 1:9 For God is my witness, whom I serve with my spirit in

the gospel of his Son, that without ceasing I make mention of you always in my prayers;

stis enim mihi est Deus cui seruio in spiritu meo in euangelio Filii eius quod sine intermissione memoriam uestri facio 1:10 semper in orationibus meis obsecrans si quo modo tan- 1:10 Making request, if by any means now at length I

might have a prosperous journey by the will of God to come unto you.

dem aliquando prosperum iter habeam in uoluntate Dei ueniendi ad uos 1:11 desidero enim uidere uos ut aliquid in- 1:11 For I long to see you, that I may impart unto you

some spiritual gift, to the end ye may be established;

pertiar gratiae uobis spiritalis ad confirmandos uos 1:12 id 1:12 That is, that I may be comforted together with you by

the mutual faith both of you and me.

est simul consolari in uobis per eam quae inuicem est fidem uestram atque meam 1:13 nolo autem uos ignorare fratres 1:13 Now I would not have you ignorant, brethren, that

oftentimes I purposed to come unto you, (but was let hitherto,) that I might have some fruit among you also, even

quia saepe proposui uenire ad uos et prohibitus sum usque as among other Gentiles. adhuc ut aliquem fructum habeam et in uobis sicut et in ceteris gentibus 1:14 Graecis ac barbaris sapientibus et insi- 1:14 I am debtor both to the Greeks, and to the Barbarians;

both to the wise, and to the unwise.

pientibus debitor sum 1:15 ita quod in me promptum est 1:15 So, as much as in me is, I am ready to preach the

gospel to you that are at Rome also.

et uobis qui Romae estis euangelizare 1:16 non enim eru- 1:16 For I am not ashamed of the gospel of Christ: for

it is the power of God unto salvation to every one that believeth; to the Jew first, and also to the Greek.

besco euangelium uirtus enim Dei est in salutem omni cre-

denti Iudaeo primum et Graeco 1:17 iustitia enim Dei in eo 1:17 For therein is the righteousness of God revealed from

faith to faith: as it is written, The just shall live by faith.

reuelatur ex fide in fidem sicut scriptum est iustus autem ex fide uiuit 1:18 reuelatur enim ira Dei de caelo super omnem 1:18 For the wrath of God is revealed from heaven against

all ungodliness and unrighteousness of men, who hold the truth in unrighteousness;

impietatem et iniustitiam hominum eorum qui ueritatem in iniustitiam detinent 1:19 quia quod notum est Dei manifes- 1:19 Because that which may be known of God is manifest

in them; for God hath shewed it unto them.

tum est in illis Deus enim illis manifestauit 1:20 inuisibilia 1:20 For the invisible things of him from the creation of

the world are clearly seen, being understood by the things that are made, even his eternal power and Godhead; so that they are without excuse:

Biblia Sacra Vulgata

enim ipsius a creatura mundi per ea quae facta sunt intellecta conspiciuntur sempiterna quoque eius uirtus et diuini-

1:21 Because that, when they knew God, they glorified tas ut sint inexcusabiles 1:21 quia cum cognouissent De-

him not as God, neither were thankful; but became vain in their imaginations, and their foolish heart was darkened.

um non sicut Deum glorificauerunt aut gratias egerunt sed euanuerunt in cogitationibus suis et obscuratum est insipiens

1:22 Professing themselves to be wise, they became fools, cor eorum 1:22 dicentes enim se esse sapientes stulti facti 1:23 And changed the glory of the uncorruptible God into sunt 1:23 et mutauerunt gloriam incorruptibilis Dei in si-

an image made like to corruptible man, and to birds, and fourfooted beasts, and creeping things.

militudinem imaginis corruptibilis hominis et uolucrum et

1:24 Wherefore God also gave them up to uncleanness quadrupedum et serpentium 1:24 propter quod tradidit illos

through the lusts of their own hearts, to dishonour their own bodies between themselves:

Deus in desideria cordis eorum in inmunditiam ut contume-

1:25 Who changed the truth of God into a lie, and wor- liis adficiant corpora sua in semet ipsis 1:25 qui commuta-

shipped and served the creature more than the Creator, who is blessed for ever. Amen.

uerunt ueritatem Dei in mendacio et coluerunt et seruierunt creaturae potius quam creatori qui est benedictus in saecula

1:26 For this cause God gave them up unto vile affections: amen 1:26 propterea tradidit illos Deus in passiones igno-

for even their women did change the natural use into that which is against nature:

miniae nam feminae eorum inmutauerunt naturalem usum

1:27 And likewise also the men, leaving the natural use of in eum usum qui est contra naturam 1:27 similiter autem et the woman, burned in their lust one toward another; men with men working that which is unseemly, and receiving in themselves that recompence of their error which was meet.

masculi relicto naturali usu feminae exarserunt in desideriis suis in inuicem masculi in masculos turpitudinem operantes et mercedem quam oportuit erroris sui in semet ipsis reci-

1:28 And even as they did not like to retain God in their pientes 1:28 et sicut non probauerunt Deum habere in no-

knowledge, God gave them over to a reprobate mind, to do those things which are not convenient;

titia tradidit eos Deus in reprobum sensum ut faciant quae

1:29 Being filled with all unrighteousness, fornication, non conueniunt 1:29 repletos omni iniquitate malitia for-

wickedness, covetousness, maliciousness; full of envy, murder, debate, deceit, malignity; whisperers,

nicatione auaritia nequitia plenos inuidia homicidio conten-

1:30 Backbiters, haters of God, despiteful, proud, boast- tione dolo malignitate susurrones 1:30 detractores Deo odi-

ers, inventors of evil things, disobedient to parents,

biles contumeliosos superbos elatos inuentores malorum pa-

1:31 Without understanding, covenantbreakers, without rentibus non oboedientes 1:31 insipientes inconpositos sine

natural affection, implacable, unmerciful:

1:32 Who knowing the judgment of God, that they which affectione absque foedere sine misericordia 1:32 qui cum

commit such things are worthy of death, not only do the same, but have pleasure in them that do them.

iustitiam Dei cognouissent non intellexerunt quoniam qui talia agunt digni sunt morte non solum ea faciunt sed et consentiunt facientibus

2:1 Therefore thou art inexcusable, O man, whosoever 2:1 propter quod inexcusabilis es o homo omnis qui iudi-

thou art that judgest: for wherein thou judgest another, thou condemnest thyself; for thou that judgest doest the

same things. cas in quo enim iudicas alterum te ipsum condemnas eadem

2:2 But we are sure that the judgment of God is according enim agis qui iudicas 2:2 scimus enim quoniam iudicium to truth against them which commit such things.

Epistola Pauli Ad Romanos

Dei est secundum ueritatem in eos qui talia agunt 2:3 ex- 2:3 And thinkest thou this, O man, that judgest them which do such things, and doest the same, that thou shalt escape the judgment of God?

istimas autem hoc o homo qui iudicas eos qui talia agunt et facis ea quia tu effugies iudicium Dei 2:4 an diuitias boni- 2:4 Or despisest thou the riches of his goodness and for-

bearance and longsuffering; not knowing that the goodness of God leadeth thee to repentance?

tatis eius et patientiae et longanimitatis contemnis ignorans

quoniam benignitas Dei ad paenitentiam te adducit 2:5 se- 2:5 But after thy hardness and impenitent heart treasurest up unto thyself wrath against the day of wrath and revelation of the righteous judgment of God;

cundum duritiam autem tuam et inpaenitens cor thesaurizas tibi iram in die irae et reuelationis iusti iudicii Dei 2:6 qui 2:6 Who will render to every man according to his deeds: reddet unicuique secundum opera eius 2:7 his quidem qui 2:7 To them who by patient continuance in well doing seek

for glory and honour and immortality, eternal life:

secundum patientiam boni operis gloriam et honorem et incorruptionem quaerentibus uitam aeternam 2:8 his autem 2:8 But unto them that are contentious, and do not obey the

truth, but obey unrighteousness, indignation and wrath,

qui ex contentione et qui non adquiescunt ueritati credunt autem iniquitati ira et indignatio 2:9 tribulatio et angustia 2:9 Tribulation and anguish, upon every soul of man that

doeth evil, of the Jew first, and also of the Gentile;

in omnem animam hominis operantis malum Iudaei primum et Graeci 2:10 gloria autem et honor et pax omni operanti 2:10 But glory, honour, and peace, to every man that wor-

keth good, to the Jew first, and also to the Gentile:

bonum Iudaeo primum et Graeco 2:11 non est enim perso- 2:11 For there is no respect of persons with God. narum acceptio apud Deum 2:12 quicumque enim sine lege 2:12 For as many as have sinned without law shall also

perish without law: and as many as have sinned in the law shall be judged by the law;

peccauerunt sine lege et peribunt et quicumque in lege peccauerunt per legem iudicabuntur 2:13 non enim auditores 2:13 (For not the hearers of the law are just before God,

but the doers of the law shall be justified.

legis iusti sunt apud Deum sed factores legis iustificabuntur 2:14 cum enim gentes quae legem non habent naturali- 2:14 For when the Gentiles, which have not the law, do by

nature the things contained in the law, these, having not the law, are a law unto themselves:

ter quae legis sunt faciunt eiusmodi legem non habentes ipsi sibi sunt lex 2:15 qui ostendunt opus legis scriptum in cor- 2:15 Which shew the work of the law written in their

hearts, their conscience also bearing witness, and their thoughts the mean while accusing or else excusing one an-

dibus suis testimonium reddente illis conscientia ipsorum et other;) inter se inuicem cogitationum accusantium aut etiam defendentium 2:16 in die cum iudicabit Deus occulta hominum 2:16 In the day when God shall judge the secrets of men

by Jesus Christ according to my gospel.

secundum euangelium meum per Iesum Christum 2:17 si 2:17 Behold, thou art called a Jew, and restest in the law,

and makest thy boast of God,

autem tu Iudaeus cognominaris et requiescis in lege et gloriaris in Deo 2:18 et nosti uoluntatem et probas utiliora in- 2:18 And knowest his will, and approvest the things that

are more excellent, being instructed out of the law;

structus per legem 2:19 confidis te ipsum ducem esse cae- 2:19 And art confident that thou thyself art a guide of the

blind, a light of them which are in darkness,

corum lumen eorum qui in tenebris sunt 2:20 eruditorem 2:20 An instructor of the foolish, a teacher of babes, which

hast the form of knowledge and of the truth in the law.

insipientium magistrum infantium habentem formam scien-

tiae et ueritatis in lege 2:21 qui ergo alium doces te ipsum 2:21 Thou therefore which teachest another, teachest thou not thyself? thou that preachest a man should not steal, dost thou steal?

non doces qui praedicas non furandum furaris 2:22 qui di-

2:22 Thou that sayest a man should not commit adultery, dost thou commit adultery? thou that abhorrest idols, dost thou commit sacrilege?

Biblia Sacra Vulgata

cis non moechandum moecharis qui abominaris idola sacri-

2:23 Thou that makest thy boast of the law, through break- legium facis 2:23 qui in lege gloriaris per praeuaricationem

ing the law dishonourest thou God?

2:24 For the name of God is blasphemed among the Gen- legis Deum inhonoras 2:24 nomen enim Dei per uos blas-

tiles through you, as it is written.

2:25 For circumcision verily profiteth, if thou keep the phematur inter gentes sicut scriptum est 2:25 circumcisio

law: but if thou be a breaker of the law, thy circumcision is made uncircumcision.

quidem prodeet si legem obserues si autem praeuaricator le-

2:26 Therefore if the uncircumcision keep the righteous- gis sis circumcisio tua praeputium facta est 2:26 si igitur

ness of the law, shall not his uncircumcision be counted for circumcision?

praeputium iustitias legis custodiat nonne praeputium illius

2:27 And shall not uncircumcision which is by nature, if it in circumcisionem reputabitur 2:27 et iudicabit quod ex na-

fulfil the law, judge thee, who by the letter and circumcision dost transgress the law?

tura est praeputium legem consummans te qui per litteram et

2:28 For he is not a Jew, which is one outwardly; neither circumcisionem praeuaricator legis es 2:28 non enim qui in

is that circumcision, which is outward in the flesh:

manifesto Iudaeus est neque quae in manifesto in carne cir-

2:29 But he is a Jew, which is one inwardly; and circumci- cumcisio 2:29 sed qui in abscondito Iudaeus et circumcisio

sion is that of the heart, in the spirit, and not in the letter; whose praise is not of men, but of God.

cordis in spiritu non littera cuius laus non ex hominibus sed ex Deo est

3:1 What advantage then hath the Jew? or what profit is 3:1 quid ergo amplius est Iudaeo aut quae utilitas circum-

there of circumcision? 3:2 Much every way: chiefly, because that unto them were cisionis 3:2 multum per omnem modum primum quidem committed the oracles of God.

3:3 For what if some did not believe? shall their unbelief quia credita sunt illis eloquia Dei 3:3 quid enim si quidam

make the faith of God without effect?

illorum non crediderunt numquid incredulitas illorum fidem

3:4 God forbid: yea, let God be true, but every man a Dei euacuabit absit 3:4 est autem Deus uerax omnis autem

liar; as it is written, That thou mightest be justified in thy sayings, and mightest overcome when thou art judged.

homo mendax sicut scriptum est ut iustificeris in sermonibus

3:5 But if our unrighteousness commend the righteousness tuis et uincas cum iudicaris 3:5 si autem iniquitas nostra iu-

of God, what shall we say? Is God unrighteous who taketh vengeance? (I speak as a man)

stitiam Dei commendat quid dicemus numquid iniquus De-

3:6 God forbid: for then how shall God judge the world? us qui infert iram secundum hominem dico 3:6 absit alio-

3:7 For if the truth of God hath more abounded through my quin quomodo iudicabit Deus mundum 3:7 si enim ueritas lie unto his glory; why yet am I also judged as a sinner?

Dei in meo mendacio abundauit in gloriam ipsius quid ad-

3:8 And not rather, (as we be slanderously reported, and huc et ego tamquam peccator iudicor 3:8 et non sicut blas-

as some affirm that we say,) Let us do evil, that good may come? whose damnation is just.

phemamur et sicut aiunt nos quidam dicere faciamus mala

3:9 What then? are we better than they? No, in no wise: ut ueniant bona quorum damnatio iusta est 3:9 quid igi-

for we have before proved both Jews and Gentiles, that they are all under sin;

tur praecellimus eos nequaquam causati enim sumus Iudaeos

3:10 As it is written, There is none righteous, no, not one: et Graecos omnes sub peccato esse 3:10 sicut scriptum est 3:11 There is none that understandeth, there is none that quia non est iustus quisquam 3:11 non est intellegens non seeketh after God.

3:12 They are all gone out of the way, they are together est requirens Deum 3:12 omnes declinauerunt simul inuti-

become unprofitable; there is none that doeth good, no, not one.

les facti sunt non est qui faciat bonum non est usque ad unum

Epistola Pauli Ad Romanos

3:13 sepulchrum patens est guttur eorum linguis suis dolose 3:13 Their throat is an open sepulchre; with their tongues

they have used deceit; the poison of asps is under their lips:

agebant uenenum aspidum sub labiis eorum 3:14 quorum

3:14 Whose mouth is full of cursing and bitterness:

os maledictione et amaritudine plenum est 3:15 ueloces pe- 3:15 Their feet are swift to shed blood: des eorum ad effundendum sanguinem 3:16 contritio et in- 3:16 Destruction and misery are in their ways: felicitas in uiis eorum 3:17 et uiam pacis non cognouerunt 3:17 And the way of peace have they not known: 3:18 non est timor Dei ante oculos eorum 3:19 scimus au- 3:18 There is no fear of God before their eyes.

3:19 Now we know that what things soever the law saith, it saith to them who are under the law: that every mouth may be stopped, and all the world may become guilty before God.

tem quoniam quaecumque lex loquitur his qui in lege sunt loquitur ut omne os obstruatur et subditus fiat omnis mundus Deo 3:20 quia ex operibus legis non iustificabitur om- 3:20 Therefore by the deeds of the law there shall no flesh

be justified in his sight: for by the law is the knowledge of sin.

nis caro coram illo per legem enim cognitio peccati 3:21

3:21 But now the righteousness of God without the law is manifested, being witnessed by the law and the prophets;

nunc autem sine lege iustitia Dei manifestata est testificata a lege et prophetis 3:22 iustitia autem Dei per fidem Iesu Ch- 3:22 Even the righteousness of God which is by faith of

Jesus Christ unto all and upon all them that believe: for there is no difference:

risti super omnes qui credunt non enim est distinctio 3:23

3:23 For all have sinned, and come short of the glory of God;

omnes enim peccauerunt et egent gloriam Dei 3:24 iustifi- 3:24 Being justified freely by his grace through the re-

demption that is in Christ Jesus:

cati gratis per gratiam ipsius per redeeptionem quae est in Christo Iesu 3:25 quem proposuit Deus propitiationem per 3:25 Whom God hath set forth to be a propitiation through

faith in his blood, to declare his righteousness for the remission of sins that are past, through the forbearance of

fidem in sanguine ipsius ad ostensionem iustitiae suae prop- God; ter remissionem praecedentium delictorum 3:26 in susten- 3:26 To declare, I say, at this time his righteousness: that

he might be just, and the justifier of him which believeth in Jesus.

tatione Dei ad ostensionem iustitiae eius in hoc tempore ut sit ipse iustus et iustificans eum qui ex fide est Iesu 3:27 ubi 3:27 Where is boasting then? It is excluded. By what law?

of works? Nay: but by the law of faith.

est ergo gloriatio exclusa est per quam legem factorum non sed per legem fidei 3:28 arbitramur enim iustificari homi- 3:28 Therefore we conclude that a man is justified by faith

without the deeds of the law.

nem per fidem sine operibus legis 3:29 an Iudaeorum Deus 3:29 Is he the God of the Jews only? is he not also of the

Gentiles? Yes, of the Gentiles also:

tantum nonne et gentium immo et gentium 3:30 quoniam 3:30 Seeing it is one God, which shall justify the circum-

cision by faith, and uncircumcision through faith.

quidem unus Deus qui iustificabit circumcisionem ex fide et praeputium per fidem 3:31 legem ergo destruimus per fi- 3:31 Do we then make void the law through faith? God

forbid: yea, we establish the law.

dem absit sed legem statuimus

4:1 quid ergo dicemus inuenisse Abraham patrem nostrum 4:1 What shall we say then that Abraham our father, as

pertaining to the flesh, hath found?

secundum carnem 4:2 si enim Abraham ex operibus iustifi- 4:2 For if Abraham were justified by works, he hath

whereof to glory; but not before God.

catus est habet gloriam sed non apud Deum 4:3 quid enim 4:3 For what saith the scripture? Abraham believed God,

and it was counted unto him for righteousness.

scriptura dicit credidit Abraham Deo et reputatum est illi ad iustitiam 4:4 ei autem qui operatur merces non inputatur 4:4 Now to him that worketh is the reward not reckoned of

grace, but of debt.

secundum gratiam sed secundum debitum 4:5 ei uero qui 4:5 But to him that worketh not, but believeth on him that justifieth the ungodly, his faith is counted for righteousness.

Biblia Sacra Vulgata

non operatur credenti autem in eum qui iustificat impium

4:6 Even as David also describeth the blessedness of the reputatur fides eius ad iustitiam 4:6 sicut et Dauid dicit be-

man, unto whom God imputeth righteousness without works,

atitudinem hominis cui Deus accepto fert iustitiam sine ope-

4:7 Saying, Blessed are they whose iniquities are forgiven, ribus 4:7 beati quorum remissae sunt iniquitates et quorum and whose sins are covered.

4:8 Blessed is the man to whom the Lord will not impute tecta sunt peccata 4:8 beatus uir cui non inputabit Domi-

sin.

4:9 Cometh this blessedness then upon the circumcision nus peccatum 4:9 beatitudo ergo haec in circumcisione an

only, or upon the uncircumcision also? for we say that faith was reckoned to Abraham for righteousness.

etiam in praeputio dicimus enim quia reputata est Abrahae

4:10 How was it then reckoned? when he was in circum- fides ad iustitiam 4:10 quomodo ergo reputata est in cir-

cision, or in uncircumcision? Not in circumcision, but in uncircumcision.

cumcisione an in praeputio non in circumcisione sed in prae-

4:11 And he received the sign of circumcision, a seal of putio 4:11 et signum accepit circumcisionis signaculum iu-

the righteousness of the faith which he had yet being uncircumcised: that he might be the father of all them that believe, though they be not circumcised; that righteousness might be imputed unto them also:

stitiae fidei quae est in praeputio ut sit pater omnium creden-

4:12 And the father of circumcision to them who are not tium per praeputium ut reputetur et illis ad iustitiam 4:12 et

of the circumcision only, but who also walk in the steps of that faith of our father Abraham, which he had being yet

uncircumcised. sit pater circumcisionis non his tantum qui sunt ex circum-

cisione sed et his qui sectantur uestigia quae est in praeputio

4:13 For the promise, that he should be the heir of the fidei patris nostri Abrahae 4:13 non enim per legem pro-

world, was not to Abraham, or to his seed, through the law, but through the righteousness of faith.

missio Abrahae aut semini eius ut heres esset mundi sed per

4:14 For if they which are of the law be heirs, faith is made iustitiam fidei 4:14 si enim qui ex lege heredes sunt exina-

void, and the promise made of none effect:

4:15 Because the law worketh wrath: for where no law is, nita est fides abolita est promissio 4:15 lex enim iram ope-

there is no transgression. 4:16 Therefore it is of faith, that it might be by grace; to ratur ubi enim non est lex nec praeuaricatio 4:16 ideo ex the end the promise might be sure to all the seed; not to that only which is of the law, but to that also which is of

the faith of Abraham; who is the father of us all, fide ut secundum gratiam ut firma sit promissio omni semini

non ei qui ex lege est solum sed et ei qui ex fide est Abra-

4:17 (As it is written, I have made thee a father of many hae qui est pater omnium nostrum 4:17 sicut scriptum est

nations,) before him whom he believed, even God, who quickeneth the dead, and calleth those things which be not

as though they were. quia patrem multarum gentium posui te ante Deum cui cre-

didit qui uiuificat mortuos et uocat quae non sunt tamquam

4:18 Who against hope believed in hope, that he might be- ea quae sunt 4:18 qui contra spem in spem credidit ut fieret

come the father of many nations, according to that which was spoken, So shall thy seed be.

pater multarum gentium secundum quod dictum est sic erit

4:19 And being not weak in faith, he considered not his semen tuum 4:19 et non infirmatus fide considerauit corpus

own body now dead, when he was about an hundred years old, neither yet the deadness of Sarah’s womb:

suum emortuum cum fere centum annorum esset et emor-

4:20 He staggered not at the promise of God through un- tuam uuluam Sarrae 4:20 in repromissione etiam Dei non

belief; but was strong in faith, giving glory to God;

haesitauit diffidentia sed confortatus est fide dans gloriam

4:21 And being fully persuaded that, what he had Deo 4:21 plenissime sciens quia quaecumque promisit po-

promised, he was able also to perform.

4:22 And therefore it was imputed to him for righteous- tens est et facere 4:22 ideo et reputatum est illi ad iustitiam

ness.

4:23 Now it was not written for his sake alone, that it was 4:23 non est autem scriptum tantum propter ipsum quia re-

imputed to him;

Epistola Pauli Ad Romanos

putatum est illi 4:24 sed et propter nos quibus reputabitur 4:24 But for us also, to whom it shall be imputed, if we

believe on him that raised up Jesus our Lord from the dead;

credentibus in eum qui suscitauit Iesum Dominum nostrum a mortuis 4:25 qui traditus est propter delicta nostra et re- 4:25 Who was delivered for our offences, and was raised

again for our justification.

surrexit propter iustificationem nostram

5:1 iustificati igitur ex fide pacem habeamus ad Deum per 5:1 Therefore being justified by faith, we have peace with

God through our Lord Jesus Christ:

Dominum nostrum Iesum Christum 5:2 per quem et acces- 5:2 By whom also we have access by faith into this grace

wherein we stand, and rejoice in hope of the glory of God.

sum habemus fide in gratiam istam in qua stamus et gloria-

mur in spe gloriae filiorum Dei 5:3 non solum autem sed 5:3 And not only so, but we glory in tribulations also:

knowing that tribulation worketh patience;

et gloriamur in tribulationibus scientes quod tribulatio patientiam operatur 5:4 patientia autem probationem probatio 5:4 And patience, experience; and experience, hope: uero spem 5:5 spes autem non confundit quia caritas Dei 5:5 And hope maketh not ashamed; because the love of

God is shed abroad in our hearts by the Holy Ghost which is given unto us.

diffusa est in cordibus nostris per Spiritum Sanctum qui datus est nobis 5:6 ut quid enim Christus cum adhuc infirmi 5:6 For when we were yet without strength, in due time

Christ died for the ungodly.

essemus secundum tempus pro impiis mortuus est 5:7 uix 5:7 For scarcely for a righteous man will one die: yet per-

adventure for a good man some would even dare to die.

enim pro iusto quis moritur nam pro bono forsitan quis et

audeat mori 5:8 commendat autem suam caritatem Deus in 5:8 But God commendeth his love toward us, in that, while

we were yet sinners, Christ died for us.

nos quoniam cum adhuc peccatores essemus 5:9 Christus 5:9 Much more then, being now justified by his blood, we

shall be saved from wrath through him.

pro nobis mortuus est multo igitur magis iustificati nunc in sanguine ipsius salui erimus ab ira per ipsum 5:10 si enim 5:10 For if, when we were enemies, we were reconciled to

God by the death of his Son, much more, being reconciled, we shall be saved by his life.

cum inimici essemus reconciliati sumus Deo per mortem Filii eius multo magis reconciliati salui erimus in uita ipsius 5:11 non solum autem sed et gloriamur in Deo per Dominum 5:11 And not only so, but we also joy in God through our

Lord Jesus Christ, by whom we have now received the atonement.

nostrum Iesum Christum per quem nunc reconciliationem accepimus 5:12 propterea sicut per unum hominem in hunc 5:12 Wherefore, as by one man sin entered into the world,

and death by sin; and so death passed upon all men, for that all have sinned:

mundum peccatum intrauit et per peccatum mors et ita in omnes homines mors pertransiit in quo omnes peccauerunt 5:13 usque ad legem enim peccatum erat in mundo peccatum 5:13 (For until the law sin was in the world: but sin is not

imputed when there is no law.

autem non inputatur cum lex non est 5:14 sed regnauit mors 5:14 Nevertheless death reigned from Adam to Moses,

even over them that had not sinned after the similitude of Adam’s transgression, who is the figure of him that was to

ab Adam usque ad Mosen etiam in eos qui non peccauerunt come. in similitudinem praeuaricationis Adae qui est forma futuri 5:15 sed non sicut delictum ita et donum si enim unius de- 5:15 But not as the offence, so also is the free gift. For if

through the offence of one many be dead, much more the grace of God, and the gift by grace, which is by one man,

licto multi mortui sunt multo magis gratia Dei et donum in Jesus Christ, hath abounded unto many. gratiam unius hominis Iesu Christi in plures abundauit 5:16 5:16 And not as it was by one that sinned, so is the gift:

for the judgment was by one to condemnation, but the free gift is of many offences unto justification.

Biblia Sacra Vulgata

et non sicut per unum peccantem ita et donum nam iudicium ex uno in condemnationem gratia autem ex multis de-

5:17 For if by one man’s offence death reigned by one; lictis in iustificationem 5:17 si enim in unius delicto mors

much more they which receive abundance of grace and of the gift of righteousness shall reign in life by one, Jesus

Christ.) regnauit per unum multo magis abundantiam gratiae et do-

nationis et iustitiae accipientes in uita regnabunt per unum

5:18 Therefore as by the offence of one judgment came Iesum Christum 5:18 igitur sicut per unius delictum in om-

upon all men to condemnation; even so by the righteousness of one the free gift came upon all men unto justifica-

tion of life. nes homines in condemnationem sic et per unius iustitiam 5:19 For as by one man’s disobedience many were made in omnes homines in iustificationem uitae 5:19 sicut enim

sinners, so by the obedience of one shall many be made righteous.

per inoboedientiam unius hominis peccatores constituti sunt multi ita et per unius oboeditionem iusti constituentur multi

5:20 Moreover the law entered, that the offence might 5:20 lex autem subintrauit ut abundaret delictum ubi autem

abound. But where sin abounded, grace did much more abound:

abundauit delictum superabundauit gratia 5:21 ut sicut reg-

5:21 That as sin hath reigned unto death, even so might grace reign through righteousness unto eternal life by Jesus Christ our Lord.

nauit peccatum in morte ita et gratia regnet per iustitiam in uitam aeternam per Iesum Christum Dominum nostrum

6:1 What shall we say then? Shall we continue in sin, that 6:1 quid ergo dicemus permanebimus in peccato ut gratia grace may abound?

6:2 God forbid. How shall we, that are dead to sin, live abundet 6:2 absit qui enim mortui sumus peccato quomodo

any longer therein?

6:3 Know ye not, that so many of us as were baptized into adhuc uiuemus in illo 6:3 an ignoratis quia quicumque ba-

Jesus Christ were baptized into his death?

ptizati sumus in Christo Iesu in morte ipsius baptizati su-

6:4 Therefore we are buried with him by baptism into mus 6:4 consepulti enim sumus cum illo per baptismum in

death: that like as Christ was raised up from the dead by the glory of the Father, even so we also should walk in

newness of life. mortem ut quomodo surrexit Christus a mortuis per gloriam

6:5 For if we have been planted together in the likeness of Patris ita et nos in nouitate uitae ambulemus 6:5 si enim his death, we shall be also in the likeness of his resurrection:

conplantati facti sumus similitudini mortis eius simul et re-

6:6 Knowing this, that our old man is crucified with him, surrectionis erimus 6:6 hoc scientes quia uetus homo nos-

that the body of sin might be destroyed, that henceforth we should not serve sin.

ter simul crucifixus est ut destruatur corpus peccati ut ultra

6:7 For he that is dead is freed from sin. non seruiamus peccato 6:7 qui enim mortuus est iustifica-6:8 Now if we be dead with Christ, we believe that we tus est a peccato 6:8 si autem mortui sumus cum Christo

shall also live with him:

6:9 Knowing that Christ being raised from the dead dieth credimus quia simul etiam uiuemus cum Christo 6:9 scien-

no more; death hath no more dominion over him.

tes quod Christus surgens ex mortuis iam non moritur mors

6:10 For in that he died, he died unto sin once: but in that illi ultra non dominabitur 6:10 quod enim mortuus est pec-

he liveth, he liveth unto God. 6:11 Likewise reckon ye also yourselves to be dead indeed cato mortuus est semel quod autem uiuit uiuit Deo 6:11 ita unto sin, but alive unto God through Jesus Christ our Lord.

et uos existimate uos mortuos quidem esse peccato uiuen-

6:12 Let not sin therefore reign in your mortal body, that tes autem Deo in Christo Iesu 6:12 non ergo regnet peccaye should obey it in the lusts thereof.

tum in uestro mortali corpore ut oboediatis concupiscentiis

Epistola Pauli Ad Romanos

eius 6:13 sed neque exhibeatis membra uestra arma iniqui- 6:13 Neither yield ye your members as instruments of un-

righteousness unto sin: but yield yourselves unto God, as those that are alive from the dead, and your members as

tatis peccato sed exhibete uos Deo tamquam ex mortuis ui- instruments of righteousness unto God. uentes et membra uestra arma iustitiae Deo 6:14 peccatum 6:14 For sin shall not have dominion over you: for ye are

not under the law, but under grace.

enim uobis non dominabitur non enim sub lege estis sed sub gratia 6:15 quid ergo peccauimus quoniam non sumus sub 6:15 What then? shall we sin, because we are not under

the law, but under grace? God forbid.

lege sed sub gratia absit 6:16 nescitis quoniam cui exhibe- 6:16 Know ye not, that to whom ye yield yourselves

servants to obey, his servants ye are to whom ye obey; whether of sin unto death, or of obedience unto righteous-

tis uos seruos ad oboediendum serui estis eius cui oboedi- ness? tis siue peccati siue oboeditionis ad iustitiam 6:17 gratias 6:17 But God be thanked, that ye were the servants of sin,

but ye have obeyed from the heart that form of doctrine which was delivered you.

autem Deo quod fuistis serui peccati oboedistis autem ex corde in eam formam doctrinae in qua traditi estis 6:18 li- 6:18 Being then made free from sin, ye became the ser-

vants of righteousness.

berati autem a peccato serui facti estis iustitiae 6:19 huma- 6:19 I speak after the manner of men because of the infir-

mity of your flesh: for as ye have yielded your members servants to uncleanness and to iniquity unto iniquity; even so now yield your members servants to righteousness unto holiness.

num dico propter infirmitatem carnis uestrae sicut enim exhibuistis membra uestra seruire inmunditiae et iniquitati ad iniquitatem ita nunc exhibete membra uestra seruire iustitiae in sanctificationem 6:20 cum enim serui essetis peccati li- 6:20 For when ye were the servants of sin, ye were free

from righteousness.

beri fuistis iustitiae 6:21 quem ergo fructum habuistis tunc 6:21 What fruit had ye then in those things whereof ye are

now ashamed? for the end of those things is death.

in quibus nunc erubescitis nam finis illorum mors est 6:22 6:22 But now being made free from sin, and become ser-

vants to God, ye have your fruit unto holiness, and the end everlasting life.

nunc uero liberati a peccato serui autem facti Deo habetis fructum uestrum in sanctificationem finem uero uitam aeternam 6:23 stipendia enim peccati mors gratia autem Dei uita 6:23 For the wages of sin is death; but the gift of God is

eternal life through Jesus Christ our Lord.

aeterna in Christo Iesu Domino nostro

7:1 an ignoratis fratres scientibus enim legem loquor quia 7:1 Know ye not, brethren, (for I speak to them that know

the law,) how that the law hath dominion over a man as long as he liveth?

lex in homine dominatur quanto tempore uiuit 7:2 nam

7:2 For the woman which hath an husband is bound by the law to her husband so long as he liveth; but if the husband be dead, she is loosed from the law of her husband.

quae sub uiro est mulier uiuente uiro alligata est legi si autem mortuus fuerit uir soluta est a lege uiri 7:3 igitur ui- 7:3 So then if, while her husband liveth, she be married to

another man, she shall be called an adulteress: but if her husband be dead, she is free from that law; so that she is

uente uiro uocabitur adultera si fuerit cum alio uiro si autem no adulteress, though she be married to another man. mortuus fuerit uir eius liberata est a lege ut non sit adultera si fuerit cum alio uiro 7:4 itaque fratres mei et uos mor- 7:4 Wherefore, my brethren, ye also are become dead to

the law by the body of Christ; that ye should be married to another, even to him who is raised from the dead, that we

tificati estis legi per corpus Christi ut sitis alterius qui ex should bring forth fruit unto God. mortuis resurrexit ut fructificaremus Deo 7:5 cum enim es- 7:5 For when we were in the flesh, the motions of sins,

which were by the law, did work in our members to bring forth fruit unto death.

semus in carne passiones peccatorum quae per legem erant operabantur in membris nostris ut fructificarent morti 7:6 7:6 But now we are delivered from the law, that being dead

wherein we were held; that we should serve in newness of spirit, and not in the oldness of the letter.

nunc autem soluti sumus a lege morientes in quo detineba-

Biblia Sacra Vulgata

mur ita ut seruiamus in nouitate spiritus et non in uetustate

7:7 What shall we say then? Is the law sin? God forbid. litterae 7:7 quid ergo dicemus lex peccatum est absit sed

Nay, I had not known sin, but by the law: for I had not known lust, except the law had said, Thou shalt not covet.

peccatum non cognoui nisi per legem nam concupiscentiam

7:8 But sin, taking occasion by the commandment, nesciebam nisi lex diceret non concupisces 7:8 occasione

wrought in me all manner of concupiscence. For without the law sin was dead.

autem accepta peccatum per mandatum operatum est in me omnem concupiscentiam sine lege enim peccatum mortuum

7:9 For I was alive without the law once: but when the erat 7:9 ego autem uiuebam sine lege aliquando sed cum

commandment came, sin revived, and I died.

7:10 And the commandment, which was ordained to life, uenisset mandatum peccatum reuixit 7:10 ego autem mor-

I found to be unto death.

tuus sum et inuentum est mihi mandatum quod erat ad uitam

7:11 For sin, taking occasion by the commandment, de- hoc esse ad mortem 7:11 nam peccatum occasione accepta

ceived me, and by it slew me.

7:12 Wherefore the law is holy, and the commandment per mandatum seduxit me et per illud occidit 7:12 itaque

holy, and just, and good.

lex quidem sancta et mandatum sanctum et iustum et bo-

7:13 Was then that which is good made death unto me? num 7:13 quod ergo bonum est mihi factum est mors absit

God forbid. But sin, that it might appear sin, working death in me by that which is good; that sin by the com-

mandment might become exceeding sinful. sed peccatum ut appareat peccatum per bonum mihi opera-

tum est mortem ut fiat supra modum peccans peccatum per

7:14 For we know that the law is spiritual: but I am carnal, mandatum 7:14 scimus enim quod lex spiritalis est ego ausold under sin. 7:15 For that which I do I allow not: for what I would, that tem carnalis sum uenundatus sub peccato 7:15 quod enim do I not; but what I hate, that do I.

operor non intellego non enim quod uolo hoc ago sed quod

7:16 If then I do that which I would not, I consent unto the odi illud facio 7:16 si autem quod nolo illud facio consen-

law that it is good.

7:17 Now then it is no more I that do it, but sin that tio legi quoniam bona 7:17 nunc autem iam non ego ope-

dwelleth in me. 7:18 For I know that in me (that is, in my flesh,) dwelleth ror illud sed quod habitat in me peccatum 7:18 scio enim no good thing: for to will is present with me; but how to perform that which is good I find not.

quia non habitat in me hoc est in carne mea bonum nam

7:19 For the good that I would I do not: but the evil which uelle adiacet mihi perficere autem bonum non inuenio 7:19

I would not, that I do.

non enim quod uolo bonum hoc facio sed quod nolo ma-

7:20 Now if I do that I would not, it is no more I that do it, lum hoc ago 7:20 si autem quod nolo illud facio non ego but sin that dwelleth in me.

7:21 I find then a law, that, when I would do good, evil is operor illud sed quod habitat in me peccatum 7:21 inuenio

present with me.

igitur legem uolenti mihi facere bonum quoniam mihi ma-

7:22 For I delight in the law of God after the inward man: lum adiacet 7:22 condelector enim legi Dei secundum in-

7:23 But I see another law in my members, warring against teriorem hominem 7:23 uideo autem aliam legem in memthe law of my mind, and bringing me into captivity to the law of sin which is in my members.

bris meis repugnantem legi mentis meae et captiuantem me

7:24 O wretched man that I am! who shall deliver me from in lege peccati quae est in membris meis 7:24 infelix ego

the body of this death?

7:25 I thank God through Jesus Christ our Lord. So then homo quis me liberabit de corpore mortis huius 7:25 gratia

with the mind I myself serve the law of God; but with the flesh the law of sin.

Dei per Iesum Christum Dominum nostrum igitur ego ipse

Epistola Pauli Ad Romanos

mente seruio legi Dei carne autem legi peccati

8:1 nihil ergo nunc damnationis est his qui sunt in Christo 8:1 There is therefore now no condemnation to them

which are in Christ Jesus, who walk not after the flesh, but after the Spirit.

Iesu qui non secundum carnem ambulant 8:2 lex enim Spi-

8:2 For the law of the Spirit of life in Christ Jesus hath made me free from the law of sin and death.

ritus uitae in Christo Iesu liberauit me a lege peccati et mortis 8:3 nam quod inpossibile erat legis in quo infirmabatur 8:3 For what the law could not do, in that it was weak

through the flesh, God sending his own Son in the likeness of sinful flesh, and for sin, condemned sin in the flesh:

per carnem Deus Filium suum mittens in similitudinem car-

nis peccati et de peccato damnauit peccatum in carne 8:4 8:4 That the righteousness of the law might be fulfilled in

us, who walk not after the flesh, but after the Spirit.

ut iustificatio legis impleretur in nobis qui non secundum

carnem ambulamus sed secundum Spiritum 8:5 qui enim 8:5 For they that are after the flesh do mind the things of the flesh; but they that are after the Spirit the things of the Spirit.

secundum carnem sunt quae carnis sunt sapiunt qui uero se-

cundum Spiritum quae sunt Spiritus sentiunt 8:6 nam pru- 8:6 For to be carnally minded is death; but to be spiritually

minded is life and peace.

dentia carnis mors prudentia autem Spiritus uita et pax 8:7 8:7 Because the carnal mind is enmity against God: for it

is not subject to the law of God, neither indeed can be.

quoniam sapientia carnis inimicitia est in Deum legi enim Dei non subicitur nec enim potest 8:8 qui autem in carne 8:8 So then they that are in the flesh cannot please God. sunt Deo placere non possunt 8:9 uos autem in carne non 8:9 But ye are not in the flesh, but in the Spirit, if so be

that the Spirit of God dwell in you. Now if any man have not the Spirit of Christ, he is none of his.

estis sed in Spiritu si tamen Spiritus Dei habitat in uobis si quis autem Spiritum Christi non habet hic non est eius 8:10 8:10 And if Christ be in you, the body is dead because of

sin; but the Spirit is life because of righteousness.

si autem Christus in uobis est corpus quidem mortuum est propter peccatum spiritus uero uita propter iustificationem 8:11 quod si Spiritus eius qui suscitauit Iesum a mortuis habi- 8:11 But if the Spirit of him that raised up Jesus from the

dead dwell in you, he that raised up Christ from the dead shall also quicken your mortal bodies by his Spirit that

tat in uobis qui suscitauit Iesum Christum a mortuis uiuifica- dwelleth in you. bit et mortalia corpora uestra propter inhabitantem Spiritum eius in uobis 8:12 ergo fratres debitores sumus non carni ut 8:12 Therefore, brethren, we are debtors, not to the flesh,

to live after the flesh.

secundum carnem uiuamus 8:13 si enim secundum carnem 8:13 For if ye live after the flesh, ye shall die: but if ye through the Spirit do mortify the deeds of the body, ye shall live.

uixeritis moriemini si autem Spiritu facta carnis mortificatis uiuetis 8:14 quicumque enim Spiritu Dei aguntur hii filii 8:14 For as many as are led by the Spirit of God, they are

the sons of God.

sunt Dei 8:15 non enim accepistis spiritum seruitutis iterum 8:15 For ye have not received the spirit of bondage again to

fear; but ye have received the Spirit of adoption, whereby we cry, Abba, Father.

in timore sed accepistis Spiritum adoptionis filiorum in quo

clamamus Abba Pater 8:16 ipse Spiritus testimonium reddit 8:16 The Spirit itself beareth witness with our spirit, that

we are the children of God:

spiritui nostro quod sumus filii Dei 8:17 si autem filii et he- 8:17 And if children, then heirs; heirs of God, and joint-

heirs with Christ; if so be that we suffer with him, that we may be also glorified together.

redes heredes quidem Dei coheredes autem Christi si tamen conpatimur ut et conglorificemur 8:18 existimo enim quod 8:18 For I reckon that the sufferings of this present time

are not worthy to be compared with the glory which shall be revealed in us.

non sunt condignae passiones huius temporis ad futuram

Biblia Sacra Vulgata

8:19 For the earnest expectation of the creature waiteth for gloriam quae reuelabitur in nobis 8:19 nam expectatio creathe manifestation of the sons of God.

8:20 For the creature was made subject to vanity, not will- turae reuelationem filiorum Dei expectat 8:20 uanitati enim

ingly, but by reason of him who hath subjected the same in hope,

creatura subiecta est non uolens sed propter eum qui subie-

8:21 Because the creature itself also shall be delivered cit in spem 8:21 quia et ipsa creatura liberabitur a seruitute

from the bondage of corruption into the glorious liberty of the children of God.

corruptionis in libertatem gloriae filiorum Dei 8:22 scimus

8:22 For we know that the whole creation groaneth and travaileth in pain together until now.

enim quod omnis creatura ingemescit et parturit usque adhuc

8:23 And not only they, but ourselves also, which have 8:23 non solum autem illa sed et nos ipsi primitias Spiritus

the firstfruits of the Spirit, even we ourselves groan within ourselves, waiting for the adoption, to wit, the redemption

of our body. habentes et ipsi intra nos gemimus adoptionem filiorum ex-8:24 For we are saved by hope: but hope that is seen is not pectantes redeeptionem corporis nostri 8:24 spe enim salui

hope: for what a man seeth, why doth he yet hope for?

facti sumus spes autem quae uidetur non est spes nam quod

8:25 But if we hope for that we see not, then do we with uidet quis quid sperat 8:25 si autem quod non uidemus spe-

patience wait for it. 8:26 Likewise the Spirit also helpeth our infirmities: for ramus per patientiam expectamus 8:26 similiter autem et we know not what we should pray for as we ought: but the Spirit itself maketh intercession for us with groanings

which cannot be uttered. Spiritus adiuuat infirmitatem nostram nam quid oremus sicut

oportet nescimus sed ipse Spiritus postulat pro nobis gemiti-

8:27 And he that searcheth the hearts knoweth what is the bus inenarrabilibus 8:27 qui autem scrutatur corda scit quid

mind of the Spirit, because he maketh intercession for the saints according to the will of God.

desideret Spiritus quia secundum Deum postulat pro sanctis

8:28 And we know that all things work together for good 8:28 scimus autem quoniam diligentibus Deum omnia coo-

to them that love God, to them who are the called according to his purpose.

perantur in bonum his qui secundum propositum uocati sunt

8:29 For whom he did foreknow, he also did predestinate sancti 8:29 nam quos praesciuit et praedestinauit confor-

to be conformed to the image of his Son, that he might be the firstborn among many brethren.

mes fieri imaginis Filii eius ut sit ipse primogenitus in mul-

8:30 Moreover whom he did predestinate, them he also tis fratribus 8:30 quos autem praedestinauit hos et uocauit

called: and whom he called, them he also justified: and whom he justified, them he also glorified.

et quos uocauit hos et iustificauit quos autem iustificauit illos

8:31 What shall we then say to these things? If God be for et glorificauit 8:31 quid ergo dicemus ad haec si Deus pro

us, who can be against us?

8:32 He that spared not his own Son, but delivered him up nobis quis contra nos 8:32 qui etiam Filio suo non peper-

for us all, how shall he not with him also freely give us all things?

cit sed pro nobis omnibus tradidit illum quomodo non etiam

8:33 Who shall lay any thing to the charge of God’s elect? cum illo omnia nobis donabit 8:33 quis accusabit aduersus

It is God that justifieth.

8:34 Who is he that condemneth? It is Christ that died, electos Dei Deus qui iustificat 8:34 quis est qui condemnet

yea rather, that is risen again, who is even at the right hand of God, who also maketh intercession for us.

Christus Iesus qui mortuus est immo qui resurrexit qui et est

8:35 Who shall separate us from the love of Christ? ad dexteram Dei qui etiam interpellat pro nobis 8:35 quis

shall tribulation, or distress, or persecution, or famine, or nakedness, or peril, or sword?

nos separabit a caritate Christi tribulatio an angustia an per-

8:36 As it is written, For thy sake we are killed all the day secutio an fames an nuditas an periculum an gladius 8:36 long; we are accounted as sheep for the slaughter.

sicut scriptum est quia propter te mortificamur tota die aes-

8:37 Nay, in all these things we are more than conquerors timati sumus ut oues occisionis 8:37 sed in his omnibus su-

through him that loved us.

Epistola Pauli Ad Romanos

peramus propter eum qui dilexit nos 8:38 certus sum enim 8:38 For I am persuaded, that neither death, nor life, nor angels, nor principalities, nor powers, nor things present, nor things to come,

quia neque mors neque uita neque angeli neque principatus neque instantia neque futura neque fortitudines 8:39 neque 8:39 Nor height, nor depth, nor any other creature, shall

be able to separate us from the love of God, which is in Christ Jesus our Lord.

altitudo neque profundum neque creatura alia poterit nos separare a caritate Dei quae est in Christo Iesu Domino nostro

9:1 ueritatem dico in Christo non mentior testimonium 9:1 I say the truth in Christ, I lie not, my conscience also

bearing me witness in the Holy Ghost,

mihi perhibente conscientia mea in Spiritu Sancto 9:2 quo- 9:2 That I have great heaviness and continual sorrow in

my heart.

niam tristitia est mihi magna et continuus dolor cordi meo 9:3 optabam enim ipse ego anathema esse a Christo pro fra- 9:3 For I could wish that myself were accursed from Christ

for my brethren, my kinsmen according to the flesh:

tribus meis qui sunt cognati mei secundum carnem 9:4 qui 9:4 Who are Israelites; to whom pertaineth the adoption, and the glory, and the covenants, and the giving of the law, and the service of God, and the promises;

sunt Israhelitae quorum adoptio est filiorum et gloria et testamenta et legislatio et obsequium et promissa 9:5 quorum 9:5 Whose are the fathers, and of whom as concerning the

flesh Christ came, who is over all, God blessed for ever. Amen.

patres et ex quibus Christus secundum carnem qui est super omnia Deus benedictus in saecula amen 9:6 non autem 9:6 Not as though the word of God hath taken none effect.

For they are not all Israel, which are of Israel:

quod exciderit uerbum Dei non enim omnes qui ex Israhel hii sunt Israhel 9:7 neque quia semen sunt Abrahae omnes 9:7 Neither, because they are the seed of Abraham, are

they all children: but, In Isaac shall thy seed be called.

filii sed in Isaac uocabitur tibi semen 9:8 id est non qui filii 9:8 That is, They which are the children of the flesh, these

are not the children of God: but the children of the promise are counted for the seed.

carnis hii filii Dei sed qui filii sunt promissionis aestimantur in semine 9:9 promissionis enim uerbum hoc est secun- 9:9 For this is the word of promise, At this time will I

come, and Sarah shall have a son.

dum hoc tempus ueniam et erit Sarrae filius 9:10 non solum 9:10 And not only this; but when Rebecca also had con-

ceived by one, even by our father Isaac;

autem sed et Rebecca ex uno concubitum habens Isaac patre nostro 9:11 cum enim nondum nati fuissent aut aliquid 9:11 (For the children being not yet born, neither having

done any good or evil, that the purpose of God according to election might stand, not of works, but of him that

egissent bonum aut malum ut secundum electionem proposi- calleth;) tum Dei maneret 9:12 non ex operibus sed ex uocante dic- 9:12 It was said unto her, The elder shall serve the

younger.

tum est ei quia maior seruiet minori 9:13 sicut scriptum est 9:13 As it is written, Jacob have I loved, but Esau have I

hated.

Iacob dilexi Esau autem odio habui 9:14 quid ergo dicemus 9:14 What shall we say then? Is there unrighteousness

with God? God forbid.

numquid iniquitas apud Deum absit 9:15 Mosi enim dicit 9:15 For he saith to Moses, I will have mercy on whom I

will have mercy, and I will have compassion on whom I will have compassion.

miserebor cuius misereor et misericordiam praestabo cuius miserebor 9:16 igitur non uolentis neque currentis sed mi- 9:16 So then it is not of him that willeth, nor of him that

runneth, but of God that sheweth mercy.

serentis Dei 9:17 dicit enim scriptura Pharaoni quia in hoc 9:17 For the scripture saith unto Pharaoh, Even for this

same purpose have I raised thee up, that I might shew my power in thee, and that my name might be declared

ipsum excitaui te ut ostendam in te uirtutem meam et ut ad- throughout all the earth. nuntietur nomen meum in uniuersa terra 9:18 ergo cuius 9:18 Therefore hath he mercy on whom he will have

mercy, and whom he will he hardeneth.

uult miseretur et quem uult indurat 9:19 dicis itaque mihi 9:19 Thou wilt say then unto me, Why doth he yet find

fault? For who hath resisted his will?

Biblia Sacra Vulgata

9:20 Nay but, O man, who art thou that repliest against quid adhuc queritur uoluntati enim eius quis resistit 9:20 o

God? Shall the thing formed say to him that formed it, Why hast thou made me thus?

homo tu quis es qui respondeas Deo numquid dicit figmen-

9:21 Hath not the potter power over the clay, of the same tum ei qui se finxit quid me fecisti sic 9:21 an non habet po-

lump to make one vessel unto honour, and another unto dishonour?

testatem figulus luti ex eadem massa facere aliud quidem uas

9:22 What if God, willing to shew his wrath, and to make in honorem aliud uero in contumeliam 9:22 quod si uolens

his power known, endured with much longsuffering the vessels of wrath fitted to destruction:

Deus ostendere iram et notam facere potentiam suam susti-

9:23 And that he might make known the riches of his glory nuit in multa patientia uasa irae aptata in interitum 9:23 ut

on the vessels of mercy, which he had afore prepared unto glory,

ostenderet diuitias gloriae suae in uasa misericordiae quae

9:24 Even us, whom he hath called, not of the Jews only, praeparauit in gloriam 9:24 quos et uocauit nos non solum

but also of the Gentiles?

9:25 As he saith also in Osee, I will call them my people, ex Iudaeis sed etiam ex gentibus 9:25 sicut in Osee dicit uo-

which were not my people; and her beloved, which was not beloved.

cabo non plebem meam plebem meam et non misericordiam

9:26 And it shall come to pass, that in the place where it consecutam misericordiam consecutam 9:26 et erit in loco

was said unto them, Ye are not my people; there shall they be called the children of the living God.

ubi dictum est eis non plebs mea uos ibi uocabuntur filii Dei

9:27 Esaias also crieth concerning Israel, Though the num- uiui 9:27 Esaias autem clamat pro Israhel si fuerit numerus ber of the children of Israel be as the sand of the sea, a remnant shall be saved:

filiorum Israhel tamquam harena maris reliquiae saluae fient

9:28 For he will finish the work, and cut it short in righ- 9:28 uerbum enim consummans et breuians in aequitate quia

teousness: because a short work will the Lord make upon the earth.

uerbum breuiatum faciet Dominus super terram 9:29 et sic-

9:29 And as Esaias said before, Except the Lord of Sabaoth had left us a seed, we had been as Sodoma, and been made like unto Gomorrha.

ut praedixit Esaias nisi Dominus Sabaoth reliquisset nobis semen sicut Sodoma facti essemus et sicut Gomorra similes

9:30 What shall we say then? That the Gentiles, which fol- fuissemus 9:30 quid ergo dicemus quod gentes quae non lowed not after righteousness, have attained to righteousness, even the righteousness which is of faith.

sectabantur iustitiam adprehenderunt iustitiam iustitiam au-

9:31 But Israel, which followed after the law of righteous- tem quae ex fide est 9:31 Israhel uero sectans legem iustiness, hath not attained to the law of righteousness.

9:32 Wherefore? Because they sought it not by faith, but tiae in legem iustitiae non peruenit 9:32 quare quia non ex

as it were by the works of the law. For they stumbled at that stumblingstone;

fide sed quasi ex operibus offenderunt in lapidem offensio-

9:33 As it is written, Behold, I lay in Sion a stumbling- nis 9:33 sicut scriptum est ecce pono in Sion lapidem of-

stone and rock of offence: and whosoever believeth on him shall not be ashamed.

fensionis et petram scandali et omnis qui credit in eum non confundetur

10:1 Brethren, my heart’s desire and prayer to God for 10:1 fratres uoluntas quidem cordis mei et obsecratio ad

Israel is, that they might be saved.

10:2 For I bear them record that they have a zeal of God, Deum fit pro illis in salutem 10:2 testimonium enim per-

but not according to knowledge.

hibeo illis quod aemulationem Dei habent sed non secun-

10:3 For they being ignorant of God’s righteousness, and dum scientiam 10:3 ignorantes enim Dei iustitiam et suam going about to establish their own righteousness, have not submitted themselves unto the righteousness of God.

quaerentes statuere iustitiae Dei non sunt subiecti 10:4 fi-

10:4 For Christ is the end of the law for righteousness to every one that believeth.

10:5 For Moses describeth the righteousness which is of nis enim legis Christus ad iustitiam omni credenti 10:5 Mo-

the law, That the man which doeth those things shall live by them.

Epistola Pauli Ad Romanos

ses enim scripsit quoniam iustitiam quae ex lege est qui fecerit homo uiuet in ea 10:6 quae autem ex fide est iustitia 10:6 But the righteousness which is of faith speaketh on

this wise, Say not in thine heart, Who shall ascend into heaven? (that is, to bring Christ down from above:)

sic dicit ne dixeris in corde tuo quis ascendit in caelum id

est Christum deducere 10:7 aut quis descendit in abyssum 10:7 Or, Who shall descend into the deep? (that is, to bring

up Christ again from the dead.)

hoc est Christum ex mortuis reuocare 10:8 sed quid dicit 10:8 But what saith it? The word is nigh thee, even in thy

mouth, and in thy heart: that is, the word of faith, which we preach;

prope est uerbum in ore tuo et in corde tuo hoc est uerbum fidei quod praedicamus 10:9 quia si confitearis in ore tuo 10:9 That if thou shalt confess with thy mouth the Lord

Jesus, and shalt believe in thine heart that God hath raised him from the dead, thou shalt be saved.

Dominum Iesum et in corde tuo credideris quod Deus il-

lum excitauit ex mortuis saluus eris 10:10 corde enim cre- 10:10 For with the heart man believeth unto righteousness;

and with the mouth confession is made unto salvation.

ditur ad iustitiam ore autem confessio fit in salutem 10:11 10:11 For the scripture saith, Whosoever believeth on him

shall not be ashamed.

dicit enim scriptura omnis qui credit in illum non confundetur 10:12 non enim est distinctio Iudaei et Graeci nam 10:12 For there is no difference between the Jew and the

Greek: for the same Lord over all is rich unto all that call upon him.

idem Dominus omnium diues in omnes qui inuocant illum 10:13 omnis enim quicumque inuocauerit nomen Domini sal- 10:13 For whosoever shall call upon the name of the Lord

shall be saved.

uus erit 10:14 quomodo ergo inuocabunt in quem non cre- 10:14 How then shall they call on him in whom they have

not believed? and how shall they believe in him of whom they have not heard? and how shall they hear without a

diderunt aut quomodo credent ei quem non audierunt quo- preacher? modo autem audient sine praedicante 10:15 quomodo uero 10:15 And how shall they preach, except they be sent? as

it is written, How beautiful are the feet of them that preach the gospel of peace, and bring glad tidings of good things!

praedicabunt nisi mittantur sicut scriptum est quam speciosi pedes euangelizantium pacem euangelizantium bona 10:16 10:16 But they have not all obeyed the gospel. For Esaias

saith, Lord, who hath believed our report?

sed non omnes oboedierunt euangelio Esaias enim dicit Domine quis credidit auditui nostro 10:17 ergo fides ex auditu 10:17 So then faith cometh by hearing, and hearing by the

word of God.

auditus autem per uerbum Christi 10:18 sed dico numquid 10:18 But I say, Have they not heard? Yes verily, their

sound went into all the earth, and their words unto the ends of the world.

non audierunt et quidem in omnem terram exiit sonus eorum et in fines orbis terrae uerba eorum 10:19 sed dico numquid 10:19 But I say, Did not Israel know? First Moses saith, I

will provoke you to jealousy by them that are no people, and by a foolish nation I will anger you.

Israhel non cognouit primus Moses dicit ego ad aemulationem uos adducam in non gentem in gentem insipientem in iram uos mittam 10:20 Esaias autem audet et dicit inuen- 10:20 But Esaias is very bold, and saith, I was found of

them that sought me not; I was made manifest unto them that asked not after me.

tus sum non quaerentibus me palam apparui his qui me non interrogabant 10:21 ad Israhel autem dicit tota die expandi 10:21 But to Israel he saith, All day long I have stretched

forth my hands unto a disobedient and gainsaying people.

manus meas ad populum non credentem et contradicentem

11:1 dico ergo numquid reppulit Deus populum suum absit 11:1 I say then, Hath God cast away his people? God for-

bid. For I also am an Israelite, of the seed of Abraham, of the tribe of Benjamin.

nam et ego Israhelita sum ex semine Abraham tribu Beniamin 11:2 non reppulit Deus plebem suam quam praesciit an 11:2 God hath not cast away his people which he

foreknew. Wot ye not what the scripture saith of Elias? how he maketh intercession to God against Israel saying,

Biblia Sacra Vulgata

nescitis in Helia quid dicit scriptura quemadmodum inter-

11:3 Lord, they have killed thy prophets, and digged down pellat Deum aduersus Israhel 11:3 Domine prophetas tuos thine altars; and I am left alone, and they seek my life.

occiderunt altaria tua suffoderunt et ego relictus sum solus et

11:4 But what saith the answer of God unto him? I have quaerunt animam meam 11:4 sed quid dicit illi responsum

reserved to myself seven thousand men, who have not bowed the knee to the image of Baal.

diuinum reliqui mihi septem milia uirorum qui non curuaue-

11:5 Even so then at this present time also there is a rem- runt genu Baal 11:5 sic ergo et in hoc tempore reliquiae se-

nant according to the election of grace.

11:6 And if by grace, then is it no more of works: other- cundum electionem gratiae factae sunt 11:6 si autem gratia

wise grace is no more grace. But if it be of works, then it is no more grace: otherwise work is no more work.

non ex operibus alioquin gratia iam non est gratia 11:7 quid

11:7 What then? Israel hath not obtained that which he seeketh for; but the election hath obtained it, and the rest were blinded.

ergo quod quaerebat Israhel hoc non est consecutus electio

11:8 (According as it is written, God hath given them the autem consecuta est ceteri uero excaecati sunt 11:8 sicut sc-

spirit of slumber, eyes that they should not see, and ears that they should not hear;) unto this day.

riptum est dedit illis Deus spiritum conpunctionis oculos ut non uideant et aures ut non audiant usque in hodiernum diem

11:9 And David saith, Let their table be made a snare, and 11:9 et Dauid dicit fiat mensa eorum in laqueum et in captio-

a trap, and a stumblingblock, and a recompence unto them:

11:10 Let their eyes be darkened, that they may not see, nem et in scandalum et in retributionem illis 11:10 obscu-

and bow down their back alway.

rentur oculi eorum ne uideant et dorsum illorum semper in-

11:11 I say then, Have they stumbled that they should fall? curua 11:11 dico ergo numquid sic offenderunt ut caderent God forbid: but rather through their fall salvation is come unto the Gentiles, for to provoke them to jealousy.

absit sed illorum delicto salus gentibus ut illos aemulentur

11:12 Now if the fall of them be the riches of the world, 11:12 quod si delictum illorum diuitiae sunt mundi et demi-

and the diminishing of them the riches of the Gentiles; how much more their fulness?

nutio eorum diuitiae gentium quanto magis plenitudo eorum

11:13 For I speak to you Gentiles, inasmuch as I am the 11:13 uobis enim dico gentibus quamdiu quidem ego sum

apostle of the Gentiles, I magnify mine office:

11:14 If by any means I may provoke to emulation them gentium apostolus ministerium meum honorificabo 11:14 si

which are my flesh, and might save some of them.

quo modo ad aemulandum prouocem carnem meam et sal-

11:15 For if the casting away of them be the reconciling uos faciam aliquos ex illis 11:15 si enim amissio eorum re-

of the world, what shall the receiving of them be, but life from the dead?

conciliatio est mundi quae adsumptio nisi uita ex mortuis

11:16 For if the firstfruit be holy, the lump is also holy: 11:16 quod si delibatio sancta est et massa et si radix sancta

and if the root be holy, so are the branches.

11:17 And if some of the branches be broken off, and thou, et rami 11:17 quod si aliqui ex ramis fracti sunt tu autem

being a wild olive tree, wert graffed in among them, and with them partakest of the root and fatness of the olive

tree; cum oleaster esses insertus es in illis et socius radicis et pin-11:18 Boast not against the branches. But if thou boast, guidinis oliuae factus es 11:18 noli gloriari aduersus ramos

thou bearest not the root, but the root thee.

11:19 Thou wilt say then, The branches were broken off, quod si gloriaris non tu radicem portas sed radix te 11:19

that I might be graffed in.

11:20 Well; because of unbelief they were broken off, and dices ergo fracti sunt rami ut ego inserar 11:20 bene pro-

thou standest by faith. Be not highminded, but fear:

pter incredulitatem fracti sunt tu autem fide stas noli altum

11:21 For if God spared not the natural branches, take sapere sed time 11:21 si enim Deus naturalibus ramis non

heed lest he also spare not thee.

11:22 Behold therefore the goodness and severity of God: pepercit ne forte nec tibi parcat 11:22 uide ergo bonitatem

on them which fell, severity; but toward thee, goodness, if thou continue in his goodness: otherwise thou also shalt be cut off.

Epistola Pauli Ad Romanos

et seueritatem Dei in eos quidem qui ceciderunt seueritatem in te autem bonitatem Dei si permanseris in bonitate alioquin et tu excideris 11:23 sed et illi si non permanserint in 11:23 And they also, if they abide not still in unbelief,

shall be graffed in: for God is able to graff them in again.

incredulitate inserentur potens est enim Deus iterum inserere illos 11:24 nam si tu ex naturali excisus es oleastro et con- 11:24 For if thou wert cut out of the olive tree which is

wild by nature, and wert graffed contrary to nature into a good olive tree: how much more shall these, which be the

tra naturam insertus es in bonam oliuam quanto magis hii natural branches, be graffed into their own olive tree?

secundum naturam inserentur suae oliuae 11:25 nolo enim 11:25 For I would not, brethren, that ye should be ignorant of this mystery, lest ye should be wise in your own conceits; that blindness in part is happened to Israel, until the

uos ignorare fratres mysterium hoc ut non sitis uobis ipsis sa- fulness of the Gentiles be come in. pientes quia caecitas ex parte contigit in Israhel donec plenitudo gentium intraret 11:26 et sic omnis Israhel saluus fieret 11:26 And so all Israel shall be saved: as it is written,

There shall come out of Sion the Deliverer, and shall turn away ungodliness from Jacob:

sicut scriptum est ueniet ex Sion qui eripiat auertet impieta-

tes ab Iacob 11:27 et hoc illis a me testamentum cum abs- 11:27 For this is my covenant unto them, when I shall take

away their sins.

tulero peccata eorum 11:28 secundum euangelium quidem 11:28 As concerning the gospel, they are enemies for your sakes: but as touching the election, they are beloved for the father’s sakes.

inimici propter uos secundum electionem autem carissimi propter patres 11:29 sine paenitentia enim sunt dona et uo- 11:29 For the gifts and calling of God are without repen-

tance.

catio Dei 11:30 sicut enim aliquando et uos non credidistis 11:30 For as ye in times past have not believed God, yet

have now obtained mercy through their unbelief:

Deo nunc autem misericordiam consecuti estis propter illorum incredulitatem 11:31 ita et isti nunc non crediderunt in 11:31 Even so have these also now not believed, that

through your mercy they also may obtain mercy.

uestram misericordiam ut et ipsi misericordiam consequantur 11:32 conclusit enim Deus omnia in incredulitatem ut 11:32 For God hath concluded them all in unbelief, that he

might have mercy upon all.

omnium misereatur 11:33 o altitudo diuitiarum sapientiae 11:33 O the depth of the riches both of the wisdom and

knowledge of God! how unsearchable are his judgments, and his ways past finding out!

et scientiae Dei quam inconprehensibilia sunt iudicia eius et inuestigabiles uiae eius 11:34 quis enim cognouit sensum 11:34 For who hath known the mind of the Lord? or who

hath been his counsellor?

Domini aut quis consiliarius eius fuit 11:35 aut quis prior 11:35 Or who hath first given to him, and it shall be rec-

ompensed unto him again?

dedit illi et retribuetur ei 11:36 quoniam ex ipso et per ip- 11:36 For of him, and through him, and to him, are all

things: to whom be glory for ever. Amen.

sum et in ipso omnia ipsi gloria in saecula amen

12:1 obsecro itaque uos fratres per misericordiam Dei ut 12:1 I beseech you therefore, brethren, by the mercies of

God, that ye present your bodies a living sacrifice, holy, acceptable unto God, which is your reasonable service.

exhibeatis corpora uestra hostiam uiuentem sanctam Deo placentem rationabile obsequium uestrum 12:2 et nolite 12:2 And be not conformed to this world: but be ye trans-

formed by the renewing of your mind, that ye may prove what is that good, and acceptable, and perfect, will of God.

conformari huic saeculo sed reformamini in nouitate sensus uestri ut probetis quae sit uoluntas Dei bona et placens et perfecta 12:3 dico enim per gratiam quae data est mihi 12:3 For I say, through the grace given unto me, to every

man that is among you, not to think of himself more highly than he ought to think; but to think soberly, according as

omnibus qui sunt inter uos non plus sapere quam oportet God hath dealt to every man the measure of faith.

sapere sed sapere ad sobrietatem unicuique sicut Deus di-

12:4 For as we have many members in one body, and all uisit mensuram fidei 12:4 sicut enim in uno corpore multa

members have not the same office:

membra habemus omnia autem membra non eundem actum

12:5 So we, being many, are one body in Christ, and every habent 12:5 ita multi unum corpus sumus in Christo sinone members one of another. 12:6 Having then gifts differing according to the grace that guli autem alter alterius membra 12:6 habentes autem dois given to us, whether prophecy, let us prophesy according to the proportion of faith;

nationes secundum gratiam quae data est nobis differentes

12:7 Or ministry, let us wait on our ministering: or he that siue prophetiam secundum rationem fidei 12:7 siue miniteacheth, on teaching; 12:8 Or he that exhorteth, on exhortation: he that giveth, sterium in ministrando siue qui docet in doctrina 12:8 qui let him do it with simplicity; he that ruleth, with diligence; he that sheweth mercy, with cheerfulness.

exhortatur in exhortando qui tribuit in simplicitate qui prae-

12:9 Let love be without dissimulation. Abhor that which est in sollicitudine qui miseretur in hilaritate 12:9 dilectio

is evil; cleave to that which is good.

12:10 Be kindly affectioned one to another with brotherly sine simulatione odientes malum adherentes bono 12:10 ca-

love; in honour preferring one another;

ritatem fraternitatis inuicem diligentes honore inuicem prae-

12:11 Not slothful in business; fervent in spirit; serving uenientes 12:11 sollicitudine non pigri spiritu feruentes Do-

the Lord; 12:12 Rejoicing in hope; patient in tribulation; continuing mino seruientes 12:12 spe gaudentes in tribulatione patieninstant in prayer;

12:13 Distributing to the necessity of saints; given to hos- tes orationi instantes 12:13 necessitatibus sanctorum com-

pitality.

12:14 Bless them which persecute you: bless, and curse municantes hospitalitatem sectantes 12:14 benedicite per-

not.

12:15 Rejoice with them that do rejoice, and weep with sequentibus benedicite et nolite maledicere 12:15 gaudere

them that weep.

12:16 Be of the same mind one toward another. Mind not cum gaudentibus flere cum flentibus 12:16 id ipsum inui-

high things, but condescend to men of low estate. Be not wise in your own conceits.

cem sentientes non alta sapientes sed humilibus consentien-

12:17 Recompense to no man evil for evil. Provide things tes nolite esse prudentes apud uosmet ipsos 12:17 nulli ma-

honest in the sight of all men.

lum pro malo reddentes prouidentes bona non tantum coram

12:18 If it be possible, as much as lieth in you, live peace- Deo sed etiam coram omnibus hominibus 12:18 si fieri poably with all men.

test quod ex uobis est cum omnibus hominibus pacem ha-

12:19 Dearly beloved, avenge not yourselves, but rather bentes 12:19 non uosmet ipsos defendentes carissimi sed give place unto wrath: for it is written, Vengeance is mine; I will repay, saith the Lord.

date locum irae scriptum est enim mihi uindictam ego ret-

12:20 Therefore if thine enemy hunger, feed him; if he ribuam dicit Dominus 12:20 sed si esurierit inimicus tuus

thirst, give him drink: for in so doing thou shalt heap coals of fire on his head.

ciba illum si sitit potum da illi hoc enim faciens carbones

12:21 Be not overcome of evil, but overcome evil with ignis congeres super caput eius 12:21 noli uinci a malo sed

good.

uince in bono malum

13:1 Let every soul be subject unto the higher powers. For 13:1 omnis anima potestatibus sublimioribus subdita sit

there is no power but of God: the powers that be are ordained of God.

non est enim potestas nisi a Deo quae autem sunt a Deo ordi-

13:2 Whosoever therefore resisteth the power, resisteth the natae sunt 13:2 itaque qui resistit potestati Dei ordinationi

ordinance of God: and they that resist shall receive to themselves damnation.

resistit qui autem resistunt ipsi sibi damnationem adquirunt

Epistola Pauli Ad Romanos

13:3 nam principes non sunt timori boni operis sed mali uis 13:3 For rulers are not a terror to good works, but to the

evil. Wilt thou then not be afraid of the power? do that which is good, and thou shalt have praise of the same:

autem non timere potestatem bonum fac et habebis laudem ex illa 13:4 Dei enim minister est tibi in bonum si autem 13:4 For he is the minister of God to thee for good. But if

thou do that which is evil, be afraid; for he beareth not the sword in vain: for he is the minister of God, a revenger to

male feceris time non enim sine causa gladium portat Dei execute wrath upon him that doeth evil. enim minister est uindex in iram ei qui malum agit 13:5 13:5 Wherefore ye must needs be subject, not only for

wrath, but also for conscience sake.

ideo necessitate subditi estote non solum propter iram sed et propter conscientiam 13:6 ideo enim et tributa praestatis 13:6 For for this cause pay ye tribute also: for they

are God’s ministers, attending continually upon this very thing.

ministri enim Dei sunt in hoc ipsum seruientes 13:7 reddite

13:7 Render therefore to all their dues: tribute to whom tribute is due; custom to whom custom; fear to whom fear; honour to whom honour.

omnibus debita cui tributum tributum cui uectigal uectigal cui timorem timorem cui honorem honorem 13:8 nemini 13:8 Owe no man any thing, but to love one another: for

he that loveth another hath fulfilled the law.

quicquam debeatis nisi ut inuicem diligatis qui enim diligit

proximum legem impleuit 13:9 nam non adulterabis non 13:9 For this, Thou shalt not commit adultery, Thou shalt not kill, Thou shalt not steal, Thou shalt not bear false witness, Thou shalt not covet; and if there be any other commandment, it is briefly comprehended in this saying, namely, Thou shalt love thy neighbour as thyself.

occides non furaberis non concupisces et si quod est aliud mandatum in hoc uerbo instauratur diliges proximum tuum

tamquam te ipsum 13:10 dilectio proximo malum non ope- 13:10 Love worketh no ill to his neighbour: therefore love is the fulfilling of the law.

ratur plenitudo ergo legis est dilectio 13:11 et hoc scien- 13:11 And that, knowing the time, that now it is high time

to awake out of sleep: for now is our salvation nearer than when we believed.

tes tempus quia hora est iam nos de somno surgere nunc enim propior est nostra salus quam cum credidimus 13:12 13:12 The night is far spent, the day is at hand: let us

therefore cast off the works of darkness, and let us put on the armour of light.

nox praecessit dies autem adpropiauit abiciamus ergo opera tenebrarum et induamur arma lucis 13:13 sicut in die ho- 13:13 Let us walk honestly, as in the day; not in rioting

and drunkenness, not in chambering and wantonness, not in strife and envying.

neste ambulemus non in comesationibus et ebrietatibus non in cubilibus et inpudicitiis non in contentione et aemulatione 13:14 sed induite Dominum Iesum Christum et carnis curam 13:14 But put ye on the Lord Jesus Christ, and make not

provision for the flesh, to fulfil the lusts thereof.

ne feceritis in desideriis

14:1 infirmum autem in fide adsumite non in disceptationi- 14:1 Him that is weak in the faith receive ye, but not to

doubtful disputations.

bus cogitationum 14:2 alius enim credit manducare omnia 14:2 For one believeth that he may eat all things: another, who is weak, eateth herbs.

qui autem infirmus est holus manducat 14:3 is qui mandu- 14:3 Let not him that eateth despise him that eateth not; and let not him which eateth not judge him that eateth: for God hath received him.

cat non manducantem non spernat et qui non manducat man-

ducantem non iudicet Deus enim illum adsumpsit 14:4 tu 14:4 Who art thou that judgest another man’s servant? to his own master he standeth or falleth. Yea, he shall be holden up: for God is able to make him stand.

quis es qui iudices alienum seruum suo domino stat aut cadit stabit autem potens est enim Deus statuere illum 14:5 nam 14:5 One man esteemeth one day above another: another

esteemeth every day alike. Let every man be fully persuaded in his own mind.

alius iudicat diem plus inter diem alius iudicat omnem diem

unusquisque in suo sensu abundet 14:6 qui sapit diem Do- 14:6 He that regardeth the day, regardeth it unto the Lord; and he that regardeth not the day, to the Lord he doth not regard it. He that eateth, eateth to the Lord, for he giveth God thanks; and he that eateth not, to the Lord he eateth not, and giveth God thanks.

Biblia Sacra Vulgata

mino sapit et qui manducat Domino manducat gratias enim agit Deo et qui non manducat Domino non manducat et gra-

14:7 For none of us liveth to himself, and no man dieth to tias agit Deo 14:7 nemo enim nostrum sibi uiuit et nemo himself.

14:8 For whether we live, we live unto the Lord; and sibi moritur 14:8 siue enim uiuimus Domino uiuimus siue

whether we die, we die unto the Lord: whether we live therefore, or die, we are the Lord’s.

morimur Domino morimur siue ergo uiuimus siue morimur

14:9 For to this end Christ both died, and rose, and revived, Domini sumus 14:9 in hoc enim Christus et mortuus est et that he might be Lord both of the dead and living.

14:10 But why dost thou judge thy brother? or why dost reuixit ut et mortuorum et uiuorum dominetur 14:10 tu au-

thou set at nought thy brother? for we shall all stand before the judgment seat of Christ.

tem quid iudicas fratrem tuum aut tu quare spernis fratrem

14:11 For it is written, As I live, saith the Lord, every knee tuum omnes enim stabimus ante tribunal Dei 14:11 scripshall bow to me, and every tongue shall confess to God.

tum est enim uiuo ego dicit Dominus quoniam mihi flec-

14:12 So then every one of us shall give account of himself tet omne genu et omnis lingua confitebitur Deo 14:12 itato God.

14:13 Let us not therefore judge one another any more: que unusquisque nostrum pro se rationem reddet Deo 14:13

but judge this rather, that no man put a stumblingblock or an occasion to fall in his brother’s way.

non ergo amplius inuicem iudicemus sed hoc iudicate magis

14:14 I know, and am persuaded by the Lord Jesus, that ne ponatis offendiculum fratri uel scandalum 14:14 scio et

there is nothing unclean of itself: but to him that esteemeth any thing to be unclean, to him it is unclean.

confido in Domino Iesu quia nihil commune per ipsum nisi

14:15 But if thy brother be grieved with thy meat, now ei qui existimat quid commune esse illi commune est 14:15

walkest thou not charitably. Destroy not him with thy meat, for whom Christ died.

si enim propter cibum frater tuus contristatur iam non secundum caritatem ambulas noli cibo tuo illum perdere pro quo

14:16 Let not then your good be evil spoken of: Christus mortuus est 14:16 non ergo blasphemetur bonum 14:17 For the kingdom of God is not meat and drink; but nostrum 14:17 non est regnum Dei esca et potus sed iustitia

righteousness, and peace, and joy in the Holy Ghost. 14:18 For he that in these things serveth Christ is accept- et pax et gaudium in Spiritu Sancto 14:18 qui enim in hoc able to God, and approved of men.

14:19 Let us therefore follow after the things which make seruit Christo placet Deo et probatus est hominibus 14:19

for peace, and things wherewith one may edify another.

itaque quae pacis sunt sectemur et quae aedificationis sunt in

14:20 For meat destroy not the work of God. All things inuicem 14:20 noli propter escam destruere opus Dei omnia

indeed are pure; but it is evil for that man who eateth with offence.

quidem munda sunt sed malum est homini qui per offendicu-

14:21 It is good neither to eat flesh, nor to drink wine, nor lum manducat 14:21 bonum est non manducare carnem et any thing whereby thy brother stumbleth, or is offended, or is made weak.

non bibere uinum neque in quo frater tuus offendit aut scan-

14:22 Hast thou faith? have it to thyself before God. dalizatur aut infirmatur 14:22 tu fidem habes penes temet

Happy is he that condemneth not himself in that thing which he alloweth.

ipsum habe coram Deo beatus qui non iudicat semet ipsum

14:23 And he that doubteth is damned if he eat, because in eo quo probat 14:23 qui autem discernit si manducauerit

he eateth not of faith: for whatsoever is not of faith is sin.

damnatus est quia non ex fide omne autem quod non ex fide peccatum est

15:1 We then that are strong ought to bear the infirmities 15:1 debemus autem nos firmiores inbecillitates infirmo-

of the weak, and not to please ourselves.

Epistola Pauli Ad Romanos

rum sustinere et non nobis placere 15:2 unusquisque ue- 15:2 Let every one of us please his neighbour for his good

to edification.

strum proximo suo placeat in bonum ad aedificationem 15:3 etenim Christus non sibi placuit sed sicut scriptum est 15:3 For even Christ pleased not himself; but, as it is writ-

ten, The reproaches of them that reproached thee fell on me.

inproperia inproperantium tibi ceciderunt super me 15:4

15:4 For whatsoever things were written aforetime were written for our learning, that we through patience and comfort of the scriptures might have hope.

quaecumque enim scripta sunt ad nostram doctrinam scripta sunt ut per patientiam et consolationem scripturarum spem habeamus 15:5 Deus autem patientiae et solacii det 15:5 Now the God of patience and consolation grant you

to be likeminded one toward another according to Christ Jesus:

uobis id ipsum sapere in alterutrum secundum Iesum Christum 15:6 ut unianimes uno ore honorificetis Deum et Pat- 15:6 That ye may with one mind and one mouth glorify

God, even the Father of our Lord Jesus Christ.

rem Domini nostri Iesu Christi 15:7 propter quod suscipite 15:7 Wherefore receive ye one another, as Christ also re-

ceived us to the glory of God.

inuicem sicut et Christus suscepit uos in honorem Dei 15:8 15:8 Now I say that Jesus Christ was a minister of the

circumcision for the truth of God, to confirm the promises made unto the fathers:

dico enim Christum Iesum ministrum fuisse circumcisionis propter ueritatem Dei ad confirmandas promissiones patrum 15:9 gentes autem super misericordiam honorare Deum sic- 15:9 And that the Gentiles might glorify God for his

mercy; as it is written, For this cause I will confess to thee among the Gentiles, and sing unto thy name.

ut scriptum est propter hoc confitebor tibi in gentibus et nomini tuo cantabo 15:10 et iterum dicit laetamini gentes cum 15:10 And again he saith, Rejoice, ye Gentiles, with his

people.

plebe eius 15:11 et iterum laudate omnes gentes Dominum 15:11 And again, Praise the Lord, all ye Gentiles; and laud

him, all ye people.

et magnificate eum omnes populi 15:12 et rursus Esaias ait 15:12 And again, Esaias saith, There shall be a root of

Jesse, and he that shall rise to reign over the Gentiles; in him shall the Gentiles trust.

erit radix Iesse et qui exsurget regere gentes in eo gentes sperabunt 15:13 Deus autem spei repleat uos omni gaudio et 15:13 Now the God of hope fill you with all joy and peace

in believing, that ye may abound in hope, through the power of the Holy Ghost.

pace in credendo ut abundetis in spe in uirtute Spiritus Sancti 15:14 certus sum autem fratres mei et ego ipse de uobis quo- 15:14 And I myself also am persuaded of you, my

brethren, that ye also are full of goodness, filled with all knowledge, able also to admonish one another.

niam et ipsi pleni estis dilectione repleti omni scientia ita ut possitis alterutrum monere 15:15 audacius autem scripsi 15:15 Nevertheless, brethren, I have written the more

boldly unto you in some sort, as putting you in mind, because of the grace that is given to me of God,

uobis fratres ex parte tamquam in memoriam uos reducens

propter gratiam quae data est mihi a Deo 15:16 ut sim mi- 15:16 That I should be the minister of Jesus Christ to the Gentiles, ministering the gospel of God, that the offering up of the Gentiles might be acceptable, being sanctified by

nister Christi Iesu in gentibus sanctificans euangelium Dei the Holy Ghost. ut fiat oblatio gentium accepta sanctificata in Spiritu Sancto 15:17 habeo igitur gloriam in Christo Iesu ad Deum 15:18 15:17 I have therefore whereof I may glory through Jesus

Christ in those things which pertain to God.

15:18 For I will not dare to speak of any of those things which Christ hath not wrought by me, to make the Gentiles obedient, by word and deed,

non enim audeo aliquid loqui eorum quae per me non effecit Christus in oboedientiam gentium uerbo et factis 15:19 _{15:19 Through mighty signs and wonders, by the power of}

the Spirit of God; so that from Jerusalem, and round about unto Illyricum, I have fully preached the gospel of Christ.

in uirtute signorum et prodiiiorum in uirtute Spiritus Sancti ita ut ab Hierusalem per circuitum usque in Illyricum rep-

Biblia Sacra Vulgata

15:20 Yea, so have I strived to preach the gospel, not leuerim euangelium Christi 15:20 sic autem hoc praedicaui

where Christ was named, lest I should build upon another man’s foundation:

euangelium non ubi nominatus est Christus ne super alienum

15:21 But as it is written, To whom he was not spoken of, fundamentum aedificarem 15:21 sed sicut scriptum est qui-

they shall see: and they that have not heard shall understand.

bus non est adnuntiatum de eo uidebunt et qui non audierunt

15:22 For which cause also I have been much hindered intellegent 15:22 propter quod et inpediebar plurimum ue-

from coming to you.

15:23 But now having no more place in these parts, and nire ad uos 15:23 nunc uero ulterius locum non habens in

having a great desire these many years to come unto you;

his regionibus cupiditatem autem habens ueniendi ad uos ex

15:24 Whensoever I take my journey into Spain, I will multis iam annis 15:24 cum in Hispaniam proficisci coe-

come to you: for I trust to see you in my journey, and to be brought on my way thitherward by you, if first I be

somewhat filled with your company. pero spero quod praeteriens uideam uos et a uobis deducar 15:25 But now I go unto Jerusalem to minister unto the illuc si uobis primum ex parte fruitus fuero 15:25 nunc igi-

saints.

15:26 For it hath pleased them of Macedonia and Achaia tur proficiscar in Hierusalem ministrare sanctis 15:26 pro-

to make a certain contribution for the poor saints which are at Jerusalem.

bauerunt enim Macedonia et Achaia conlationem aliquam

15:27 It hath pleased them verily; and their debtors they facere in pauperes sanctorum qui sunt in Hierusalem 15:27

are. For if the Gentiles have been made partakers of their spiritual things, their duty is also to minister unto them in

carnal things. placuit enim eis et debitores sunt eorum nam si spiritalium

eorum participes facti sunt gentiles debent et in carnalibus

15:28 When therefore I have performed this, and have ministrare eis 15:28 hoc igitur cum consummauero et adsi-

sealed to them this fruit, I will come by you into Spain.

gnauero eis fructum hunc proficiscar per uos in Hispaniam

15:29 And I am sure that, when I come unto you, I shall 15:29 scio autem quoniam ueniens ad uos in abundantia be-

come in the fulness of the blessing of the gospel of Christ.

15:30 Now I beseech you, brethren, for the Lord Jesus nedictionis Christi ueniam 15:30 obsecro igitur uos fratres

Christ’s sake, and for the love of the Spirit, that ye strive together with me in your prayers to God for me;

per Dominum nostrum Iesum Christum et per caritatem Spi-

15:31 That I may be delivered from them that do not be- ritus ut adiuuetis me in orationibus pro me ad Deum 15:31

lieve in Judaea; and that my service which I have for Jerusalem may be accepted of the saints;

ut liberer ab infidelibus qui sunt in Iudaea et obsequii mei

15:32 That I may come unto you with joy by the will of oblatio accepta fiat in Hierosolyma sanctis 15:32 ut ueniam

God, and may with you be refreshed.

ad uos in gaudio per uoluntatem Dei et refrigerer uobiscum

15:33 Now the God of peace be with you all. Amen. 15:33 Deus autem pacis sit cum omnibus uobis amen

16:1 I commend unto you Phebe our sister, which is a ser- 16:1 commendo autem uobis Phoebem sororem nostram vant of the church which is at Cenchrea:

16:2 That ye receive her in the Lord, as becometh saints, quae est in ministerio ecclesiae quae est Cenchris 16:2 ut

and that ye assist her in whatsoever business she hath need of you: for she hath been a succourer of many, and of

myself also. eam suscipiatis in Domino digne sanctis et adsistatis ei in

quocumque negotio uestri indiguerit etenim ipsa quoque ad-

16:3 Greet Priscilla and Aquila my helpers in Christ Jesus: stitit multis et mihi ipsi 16:3 salutate Priscam et Aquilam

16:4 Who have for my life laid down their own necks: unto adiutores meos in Christo Iesu 16:4 qui pro anima mea suas

whom not only I give thanks, but also all the churches of the Gentiles.

ceruices subposuerunt quibus non solus ego gratias ago sed

16:5 Likewise greet the church that is in their house. et cunctae ecclesiae gentium 16:5 et domesticam eorum ec-

Salute my well-beloved Epaenetus, who is the firstfruits of Achaia unto Christ.

Epistola Pauli Ad Romanos

clesiam salutate Ephaenetum dilectum mihi qui est primitiuus Asiae in Christo 16:6 salutate Mariam quae multum 16:6 Greet Mary, who bestowed much labour on us. laborauit in uobis 16:7 salutate Andronicum et Iuniam co- 16:7 Salute Andronicus and Junia, my kinsmen, and my

fellow-prisoners, who are of note among the apostles, who also were in Christ before me.

gnatos et concaptiuos meos qui sunt nobiles in apostolis qui et ante me fuerunt in Christo 16:8 salutate Ampliatum di- 16:8 Greet Amplias my beloved in the Lord. lectissimum mihi in Domino 16:9 salutate Vrbanum adiu- 16:9 Salute Urbane, our helper in Christ, and Stachys my

beloved.

torem nostrum in Christo et Stachyn dilectum meum 16:10 16:10 Salute Apelles approved in Christ. Salute them

which are of Aristobulus’ household.

salutate Apellen probum in Christo 16:11 salutate eos qui 16:11 Salute Herodion my kinsman. Greet them that be of

the household of Narcissus, which are in the Lord.

sunt ex Aristoboli salutate Herodionem cognatum meum salutate eos qui sunt ex Narcissi qui sunt in Domino 16:12 16:12 Salute Tryphena and Tryphosa, who labour in the

Lord. Salute the beloved Persis, which laboured much in the Lord.

salutate Tryfenam et Tryfosam quae laborant in Domino salutate Persidam carissimam quae multum laborauit in Do-

mino 16:13 salutate Rufum electum in Domino et matrem 16:13 Salute Rufus chosen in the Lord, and his mother and

mine.

eius et meam 16:14 salutate Asyncritum Flegonta Hermen 16:14 Salute Asyncritus, Phlegon, Hermas, Patrobas, Her-

mes, and the brethren which are with them.

Patrobam Hermam et qui cum eis sunt fratres 16:15 salu- 16:15 Salute Philologus, and Julia, Nereus, and his sister,

and Olympas, and all the saints which are with them.

tate Filologum et Iuliam Nereum et sororem eius et Olympiadem et omnes qui cum eis sunt sanctos 16:16 salutate in- 16:16 Salute one another with an holy kiss. The churches

of Christ salute you.

uicem in osculo sancto salutant uos omnes ecclesiae Christi 16:17 rogo autem uos fratres ut obseruetis eos qui dissen- 16:17 Now I beseech you, brethren, mark them which

cause divisions and offences contrary to the doctrine which ye have learned; and avoid them.

siones et offendicula praeter doctrinam quam uos didicistis faciunt et declinate ab illis 16:18 huiusmodi enim Christo 16:18 For they that are such serve not our Lord Jesus

Christ, but their own belly; and by good words and fair speeches deceive the hearts of the simple.

Domino nostro non seruiunt sed suo uentri et per dulces sermones et benedictiones seducunt corda innocentium 16:19 16:19 For your obedience is come abroad unto all men. I

am glad therefore on your behalf: but yet I would have you wise unto that which is good, and simple concerning

uestra enim oboedientia in omnem locum diuulgata est gau- evil. deo igitur in uobis sed uolo uos sapientes esse in bono et simplices in malo 16:20 Deus autem pacis conteret Satanan sub 16:20 And the God of peace shall bruise Satan under your

feet shortly. The grace of our Lord Jesus Christ be with you. Amen.

pedibus uestris uelociter gratia Domini nostri Iesu Christi uobiscum 16:21 salutat uos Timotheus adiutor meus et Lu- 16:21 Timotheus my workfellow, and Lucius, and Jason,

and Sosipater, my kinsmen, salute you.

cius et Iason et Sosipater cognati mei 16:22 saluto uos ego 16:22 I Tertius, who wrote this epistle, salute you in the

Lord.

Tertius qui scripsi epistulam in Domino 16:23 salutat uos 16:23 Gaius mine host, and of the whole church, saluteth

you. Erastus the chamberlain of the city saluteth you, and Quartus a brother.

Gaius hospes meus et uniuersae ecclesiae salutat uos Erastus arcarius ciuitatis et Quartus frater 16:24 16:25 ei au- 16:24 The grace of our Lord Jesus Christ be with you all.

Amen.

16:25 Now to him that is of power to stablish you according to my gospel, and the preaching of Jesus Christ, according to the revelation of the mystery, which was kept secret since the world began,

tem qui potens est uos confirmare iuxta euangelium meum

Biblia Sacra Vulgata

et praedicationem Iesu Christi secundum reuelationem my-

16:26 But now is made manifest, and by the scriptures of sterii temporibus aeternis taciti 16:26 quod nunc patefac-

the prophets, according to the commandment of the everlasting God, made known to all nations for the obedience

of faith: tum est per scripturas prophetarum secundum praeceptum

aeterni Dei ad oboeditionem fidei in cunctis gentibus cog-

16:27 To God only wise, be glory through Jesus Christ for nito 16:27 solo sapienti Deo per Iesum Christum cui honor

ever. Amen.

in saecula saeculorum amen

Epistola Pauli Ad Corinthios I

1:1 Paulus uocatus apostolus Christi Iesu per uoluntatem 1:1 Paul called to be an apostle of Jesus Christ through the

will of God, and Sosthenes our brother,

Dei et Sosthenes frater 1:2 ecclesiae Dei quae est Corinthi 1:2 Unto the church of God which is at Corinth, to them

that are sanctified in Christ Jesus, called to be saints, with all that in every place call upon the name of Jesus Christ

sanctificatis in Christo Iesu uocatis sanctis cum omnibus qui our Lord, both their’s and our’s: inuocant nomen Domini nostri Iesu Christi in omni loco ipsorum et nostro 1:3 gratia uobis et pax a Deo Patre nostro 1:3 Grace be unto you, and peace, from God our Father,

and from the Lord Jesus Christ.

et Domino Iesu Christo 1:4 gratias ago Deo meo semper 1:4 I thank my God always on your behalf, for the grace

of God which is given you by Jesus Christ;

pro uobis in gratia Dei quae data est uobis in Christo Iesu 1:5 quia in omnibus diuites facti estis in illo in omni uerbo et 1:5 That in every thing ye are enriched by him, in all ut-

terance, and in all knowledge;

in omni scientia 1:6 sicut testimonium Christi confirmatum 1:6 Even as the testimony of Christ was confirmed in you: est in uobis 1:7 ita ut nihil uobis desit in ulla gratia expec- 1:7 So that ye come behind in no gift; waiting for the com-

ing of our Lord Jesus Christ:

tantibus reuelationem Domini nostri Iesu Christi 1:8 qui et 1:8 Who shall also confirm you unto the end, that ye may

be blameless in the day of our Lord Jesus Christ.

confirmabit uos usque ad finem sine crimine in die aduentus Domini nostri Iesu Christi 1:9 fidelis Deus per quem uocati 1:9 God is faithful, by whom ye were called unto the fel-

lowship of his Son Jesus Christ our Lord.

estis in societatem Filii eius Iesu Christi Domini nostri 1:10 1:10 Now I beseech you, brethren, by the name of our

Lord Jesus Christ, that ye all speak the same thing, and that there be no divisions among you; but that ye be perfectly joined together in the same mind and in the same judgment.

obsecro autem uos fratres per nomen Domini nostri Iesu Christi ut id ipsum dicatis omnes et non sint in uobis scismata sitis autem perfecti in eodem sensu et in eadem sententia 1:11 significatum est enim mihi de uobis fratres mei ab his 1:11 For it hath been declared unto me of you, my

brethren, by them which are of the house of Chloe, that there are contentions among you.

qui sunt Chloes quia contentiones inter uos sunt 1:12 hoc

1:12 Now this I say, that every one of you saith, I am of Paul; and I of Apollos; and I of Cephas; and I of Christ.

autem dico quod unusquisque uestrum dicit ego quidem sum Pauli ego autem Apollo ego uero Cephae ego autem Christi 1:13 diuisus est Christus numquid Paulus crucifixus est pro 1:13 Is Christ divided? was Paul crucified for you? or

were ye baptized in the name of Paul?

uobis aut in nomine Pauli baptizati estis 1:14 gratias ago 1:14 I thank God that I baptized none of you, but Crispus

and Gaius;

Deo quod neminem uestrum baptizaui nisi Crispum et Gaium 1:15 ne quis dicat quod in nomine meo baptizati sitis 1:15 Lest any should say that I had baptized in mine own

name.

1:16 baptizaui autem et Stephanae domum ceterum nescio 1:16 And I baptized also the household of Stephanas: be-

sides, I know not whether I baptized any other.

si quem alium baptizauerim 1:17 non enim misit me Ch- 1:17 For Christ sent me not to baptize, but to preach the gospel: not with wisdom of words, lest the cross of Christ should be made of none effect.

ristus baptizare sed euangelizare non in sapientia uerbi ut non euacuetur crux Christi 1:18 uerbum enim crucis per- 1:18 For the preaching of the cross is to them that perish

foolishness; but unto us which are saved it is the power of God.

euntibus quidem stultitia est his autem qui salui fiunt id est

nobis uirtus Dei est 1:19 scriptum est enim perdam sapien- 1:19 For it is written, I will destroy the wisdom of the wise,

and will bring to nothing the understanding of the prudent.

tiam sapientium et prudentiam prudentium reprobabo 1:20 1:20 Where is the wise? where is the scribe? where is

the disputer of this world? hath not God made foolish the wisdom of this world?

ubi sapiens ubi scriba ubi conquisitor huius saeculi nonne

Biblia Sacra Vulgata

1:21 For after that in the wisdom of God the world by stultam fecit Deus sapientiam huius mundi 1:21 nam quia

wisdom knew not God, it pleased God by the foolishness of preaching to save them that believe.

in Dei sapientia non cognouit mundus per sapientiam Deum placuit Deo per stultitiam praedicationis saluos facere

1:22 For the Jews require a sign, and the Greeks seek after credentes 1:22 quoniam et Iudaei signa petunt et Graeci sa-

wisdom: 1:23 But we preach Christ crucified, unto the Jews a stum- pientiam quaerunt 1:23 nos autem praedicamus Christum blingblock, and unto the Greeks foolishness;

crucifixum Iudaeis quidem scandalum gentibus autem stul-

1:24 But unto them which are called, both Jews and titiam 1:24 ipsis autem uocatis Iudaeis atque Graecis Chris-

Greeks, Christ the power of God, and the wisdom of God.

1:25 Because the foolishness of God is wiser than men; tum Dei uirtutem et Dei sapientiam 1:25 quia quod stultum

and the weakness of God is stronger than men.

est Dei sapientius est hominibus et quod infirmum est Dei

1:26 For ye see your calling, brethren, how that not many fortius est hominibus 1:26 uidete enim uocationem uestram

wise men after the flesh, not many mighty, not many noble, are called:

fratres quia non multi sapientes secundum carnem non multi

1:27 But God hath chosen the foolish things of the world potentes non multi nobiles 1:27 sed quae stulta sunt mundi

to confound the wise; and God hath chosen the weak things of the world to confound the things which are

mighty; elegit Deus ut confundat sapientes et infirma mundi elegit 1:28 And base things of the world, and things which are Deus ut confundat fortia 1:28 et ignobilia mundi et con-

despised, hath God chosen, yea, and things which are not, to bring to nought things that are:

temptibilia elegit Deus et quae non sunt ut ea quae sunt de-

1:29 That no flesh should glory in his presence. strueret 1:29 ut non glorietur omnis caro in conspectu eius 1:30 But of him are ye in Christ Jesus, who of God is made 1:30 ex ipso autem uos estis in Christo Iesu qui factus est

unto us wisdom, and righteousness, and sanctification, and redemption:

sapientia nobis a Deo et iustitia et sanctificatio et redeeptio

1:31 That, according as it is written, He that glorieth, let 1:31 ut quemadmodum scriptum est qui gloriatur in Domino him glory in the Lord.

glorietur

2:1 And I, brethren, when I came to you, came not with 2:1 et ego cum uenissem ad uos fratres ueni non per su-

excellency of speech or of wisdom, declaring unto you the testimony of God.

blimitatem sermonis aut sapientiae adnuntians uobis testi-

2:2 For I determined not to know any thing among you, monium Christi 2:2 non enim iudicaui scire me aliquid in-

save Jesus Christ, and him crucified.

2:3 And I was with you in weakness, and in fear, and in ter uos nisi Iesum Christum et hunc crucifixum 2:3 et ego

much trembling.

2:4 And my speech and my preaching was not with entic- in infirmitate et timore et tremore multo fui apud uos 2:4

ing words of man’s wisdom, but in demonstration of the Spirit and of power:

et sermo meus et praedicatio mea non in persuasibilibus

2:5 That your faith should not stand in the wisdom of men, sapientiae uerbis sed in ostensione Spiritus et uirtutis 2:5

but in the power of God.

ut fides uestra non sit in sapientia hominum sed in uirtute

2:6 Howbeit we speak wisdom among them that are per- Dei 2:6 sapientiam autem loquimur inter perfectos sapien-

fect: yet not the wisdom of this world, nor of the princes of this world, that come to nought:

tiam uero non huius saeculi neque principum huius saeculi

2:7 But we speak the wisdom of God in a mystery, even qui destruuntur 2:7 sed loquimur Dei sapientiam in myste-

the hidden wisdom, which God ordained before the world unto our glory:

rio quae abscondita est quam praedestinauit Deus ante sae-

2:8 Which none of the princes of this world knew: for had cula in gloriam nostram 2:8 quam nemo principum huius they known it, they would not have crucified the Lord of glory.

Epistola Pauli Ad Corinthios I

saeculi cognouit si enim cognouissent numquam Dominum gloriae crucifixissent 2:9 sed sicut scriptum est quod oculus 2:9 But as it is written, Eye hath not seen, nor ear heard,

neither have entered into the heart of man, the things which God hath prepared for them that love him.

non uidit nec auris audiuit nec in cor hominis ascendit quae praeparauit Deus his qui diligunt illum 2:10 nobis autem 2:10 But God hath revealed them unto us by his Spirit: for

the Spirit searcheth all things, yea, the deep things of God.

reuelauit Deus per Spiritum suum Spiritus enim omnia scrutatur etiam profunda Dei 2:11 quis enim scit hominum quae 2:11 For what man knoweth the things of a man, save the

spirit of man which is in him? even so the things of God knoweth no man, but the Spirit of God.

sint hominis nisi spiritus hominis qui in ipso est ita et quae

Dei sunt nemo cognouit nisi Spiritus Dei 2:12 nos autem 2:12 Now we have received, not the spirit of the world, but the spirit which is of God; that we might know the things that are freely given to us of God.

non spiritum mundi accepimus sed Spiritum qui ex Deo est ut sciamus quae a Deo donata sunt nobis 2:13 quae et loqui- 2:13 Which things also we speak, not in the words which

man’s wisdom teacheth, but which the Holy Ghost teacheth; comparing spiritual things with spiritual.

mur non in doctis humanae sapientiae uerbis sed in doctrina Spiritus spiritalibus spiritalia conparantes 2:14 animalis au- 2:14 But the natural man receiveth not the things of the

Spirit of God: for they are foolishness unto him: neither can he know them, because they are spiritually discerned.

tem homo non percipit ea quae sunt Spiritus Dei stultitia est enim illi et non potest intellegere quia spiritaliter examinatur 2:15 spiritalis autem iudicat omnia et ipse a nemine iudicatur 2:15 But he that is spiritual judgeth all things, yet he him-

self is judged of no man.

2:16 quis enim cognouit sensum Domini qui instruat eum nos 2:16 For who hath known the mind of the Lord, that he

may instruct him? But we have the mind of Christ.

autem sensum Christi habemus

3:1 et ego fratres non potui uobis loqui quasi spiritalibus 3:1 And I, brethren, could not speak unto you as unto spir-

itual, but as unto carnal, even as unto babes in Christ.

sed quasi carnalibus tamquam paruulis in Christo 3:2 lac 3:2 I have fed you with milk, and not with meat: for hith-

erto ye were not able to bear it, neither yet now are ye able.

uobis potum dedi non escam nondum enim poteratis sed ne nunc quidem potestis adhuc enim estis carnales 3:3 cum 3:3 For ye are yet carnal: for whereas there is among you

envying, and strife, and divisions, are ye not carnal, and walk as men?

enim sit inter uos zelus et contentio nonne carnales estis

et secundum hominem ambulatis 3:4 cum enim quis dicit 3:4 For while one saith, I am of Paul; and another, I am of

Apollos; are ye not carnal?

ego quidem sum Pauli alius autem ego Apollo nonne homines estis quid igitur est Apollo quid uero Paulus 3:5 mi- 3:5 Who then is Paul, and who is Apollos, but ministers by

whom ye believed, even as the Lord gave to every man?

nistri eius cui credidistis et unicuique sicut Dominus dedit 3:6 ego plantaui Apollo rigauit sed Deus incrementum de- 3:6 I have planted, Apollos watered; but God gave the in-

crease.

dit 3:7 itaque neque qui plantat est aliquid neque qui ri- 3:7 So then neither is he that planteth any thing, neither he

that watereth; but God that giveth the increase.

gat sed qui incrementum dat Deus 3:8 qui plantat autem 3:8 Now he that planteth and he that watereth are one: and every man shall receive his own reward according to his own labour.

et qui rigat unum sunt unusquisque autem propriam merce-

dem accipiet secundum suum laborem 3:9 Dei enim sumus 3:9 For we are labourers together with God: ye are God’s

husbandry, ye are God’s building.

adiutores Dei agricultura estis Dei aedificatio estis 3:10 se- 3:10 According to the grace of God which is given unto

me, as a wise masterbuilder, I have laid the foundation, and another buildeth thereon. But let every man take heed

cundum gratiam Dei quae data est mihi ut sapiens architec- how he buildeth thereupon.

Biblia Sacra Vulgata

tus fundamentum posui alius autem superaedificat unusquis-

3:11 For other foundation can no man lay than that is laid, que autem uideat quomodo superaedificet 3:11 fundamenwhich is Jesus Christ.

tum enim aliud nemo potest ponere praeter id quod positum

3:12 Now if any man build upon this foundation gold, sil- est qui est Christus Iesus 3:12 si quis autem superaedificat

ver, precious stones, wood, hay, stubble;

supra fundamentum hoc aurum argentum lapides pretiosos

3:13 Every man’s work shall be made manifest: for the ligna faenum stipulam 3:13 uniuscuiusque opus manifes-

day shall declare it, because it shall be revealed by fire; and the fire shall try every man’s work of what sort it is.

tum erit dies enim declarabit quia in igne reuelabitur et uniu-

3:14 If any man’s work abide which he hath built there- scuiusque opus quale sit ignis probabit 3:14 si cuius opus

upon, he shall receive a reward.

3:15 If any man’s work shall be burned, he shall suffer manserit quod superaedificauit mercedem accipiet 3:15 si

loss: but he himself shall be saved; yet so as by fire.

cuius opus arserit detrimentum patietur ipse autem saluus

3:16 Know ye not that ye are the temple of God, and that erit sic tamen quasi per ignem 3:16 nescitis quia templum

the Spirit of God dwelleth in you?

3:17 If any man defile the temple of God, him shall God Dei estis et Spiritus Dei habitat in uobis 3:17 si quis au-

destroy; for the temple of God is holy, which temple ye are.

tem templum Dei uiolauerit disperdet illum Deus templum

3:18 Let no man deceive himself. If any man among you enim Dei sanctum est quod estis uos 3:18 nemo se seducat

seemeth to be wise in this world, let him become a fool, that he may be wise.

si quis uidetur inter uos sapiens esse in hoc saeculo stultus

3:19 For the wisdom of this world is foolishness with God. fiat ut sit sapiens 3:19 sapientia enim huius mundi stultitia

For it is written, He taketh the wise in their own craftiness.

est apud Deum scriptum est enim conprehendam sapientes

3:20 And again, The Lord knoweth the thoughts of the in astutia eorum 3:20 et iterum Dominus nouit cogitationes

wise, that they are vain.

3:21 Therefore let no man glory in men. For all things are sapientium quoniam uanae sunt 3:21 itaque nemo glorietur

your’s;

3:22 Whether Paul, or Apollos, or Cephas, or the world, in hominibus omnia enim uestra sunt 3:22 siue Paulus siue

or life, or death, or things present, or things to come; all are your’s;

Apollo siue Cephas siue mundus siue uita siue mors siue

3:23 And ye are Christ’s; and Christ is God’s. praesentia siue futura omnia enim uestra sunt 3:23 uos autem Christi Christus autem Dei

4:1 Let a man so account of us, as of the ministers of 4:1 sic nos existimet homo ut ministros Christi et dispen-

Christ, and stewards of the mysteries of God.

4:2 Moreover it is required in stewards, that a man be satores mysteriorum Dei 4:2 hic iam quaeritur inter dispen-

found faithful.

4:3 But with me it is a very small thing that I should be satores ut fidelis quis inueniatur 4:3 mihi autem pro mi-

judged of you, or of man’s judgment: yea, I judge not mine own self.

nimo est ut a uobis iudicer aut ab humano die sed neque me

4:4 For I know nothing by myself; yet am I not hereby ipsum iudico 4:4 nihil enim mihi conscius sum sed non in

justified: but he that judgeth me is the Lord.

4:5 Therefore judge nothing before the time, until the hoc iustificatus sum qui autem iudicat me Dominus est 4:5

Lord come, who both will bring to light the hidden things of darkness, and will make manifest the counsels of the

hearts: and then shall every man have praise of God. itaque nolite ante tempus iudicare quoadusque ueniat Domi-

nus qui et inluminabit abscondita tenebrarum et manifestabit

4:6 And these things, brethren, I have in a figure trans- consilia cordium et tunc laus erit unicuique a Deo 4:6 haec

ferred to myself and to Apollos for your sakes; that ye might learn in us not to think of men above that which is written, that no one of you be puffed up for one against another.

autem fratres transfiguraui in me et Apollo propter uos ut in

Epistola Pauli Ad Corinthios I

nobis discatis ne supra quam scriptum est unus aduersus alterum infletur pro alio 4:7 quis enim te discernit quid autem 4:7 For who maketh thee to differ from another? and what

hast thou that thou didst not receive? now if thou didst receive it, why dost thou glory, as if thou hadst not received

habes quod non accepisti si autem accepisti quid gloriaris it? quasi non acceperis 4:8 iam saturati estis iam diuites facti 4:8 Now ye are full, now ye are rich, ye have reigned as

kings without us: and I would to God ye did reign, that we also might reign with you.

estis sine nobis regnastis et utinam regnaretis ut et nos uobi-

scum regnaremus 4:9 puto enim Deus nos apostolos nouis- 4:9 For I think that God hath set forth us the apostles last, as it were appointed to death: for we are made a spectacle unto the world, and to angels, and to men.

simos ostendit tamquam morti destinatos quia spectaculum facti sumus mundo et angelis et hominibus 4:10 nos stulti 4:10 We are fools for Christ’s sake, but ye are wise in

Christ; we are weak, but ye are strong; ye are honourable, but we are despised.

propter Christum uos autem prudentes in Christo nos infirmi uos autem fortes uos nobiles nos autem ignobiles 4:11 us- 4:11 Even unto this present hour we both hunger, and

thirst, and are naked, and are buffeted, and have no certain dwellingplace;

que in hanc horam et esurimus et sitimus et nudi sumus et colaphis caedimur et instabiles sumus 4:12 et laboramus ope- 4:12 And labour, working with our own hands: being re-

viled, we bless; being persecuted, we suffer it:

rantes manibus nostris maledicimur et benedicimus persecutionem patimur et sustinemus 4:13 blasphemamur et obse- 4:13 Being defamed, we intreat: we are made as the filth

of the world, and are the offscouring of all things unto this day.

cramus tamquam purgamenta huius mundi facti sumus omnium peripsima usque adhuc 4:14 non ut confundam uos 4:14 I write not these things to shame you, but as my

beloved sons I warn you.

haec scribo sed ut filios meos carissimos moneo 4:15 nam 4:15 For though ye have ten thousand instructers in Christ, yet have ye not many fathers: for in Christ Jesus I have begotten you through the gospel.

si decem milia pedagogorum habeatis in Christo sed non multos patres nam in Christo Iesu per euangelium ego uos genui 4:16 rogo ergo uos imitatores mei estote 4:17 ideo 4:16 Wherefore I beseech you, be ye followers of me.

4:17 For this cause have I sent unto you Timotheus, who is my beloved son, and faithful in the Lord, who shall bring you into remembrance of my ways which be in Christ, as I teach every where in every church.

misi ad uos Timotheum qui est filius meus carissimus et fidelis in Domino qui uos commonefaciat uias meas quae sunt in Christo sicut ubique in omni ecclesia doceo 4:18 tam- 4:18 Now some are puffed up, as though I would not come

to you.

quam non uenturus sim ad uos sic inflati sunt quidam 4:19 4:19 But I will come to you shortly, if the Lord will, and

will know, not the speech of them which are puffed up, but the power.

ueniam autem cito ad uos si Dominus uoluerit et cognoscam non sermonem eorum qui inflati sunt sed uirtutem 4:20 non 4:20 For the kingdom of God is not in word, but in power. enim in sermone est regnum Dei sed in uirtute 4:21 quid 4:21 What will ye? shall I come unto you with a rod, or in

love, and in the spirit of meekness?

uultis in uirga ueniam ad uos an in caritate et spiritu mansuetudinis

5:1 omnino auditur inter uos fornicatio et talis fornicatio 5:1 It is reported commonly that there is fornication

among you, and such fornication as is not so much as named among the Gentiles, that one should have his fa-

qualis nec inter gentes ita ut uxorem patris aliquis habeat ther’s wife. 5:2 et uos inflati estis et non magis luctum habuistis ut tol- 5:2 And ye are puffed up, and have not rather mourned,

that he that hath done this deed might be taken away from among you.

latur de medio uestrum qui hoc opus fecit 5:3 ego quidem

5:3 For I verily, as absent in body, but present in spirit, have judged already, as though I were present, concerning him that hath so done this deed,

Biblia Sacra Vulgata

absens corpore praesens autem spiritu iam iudicaui ut prae-

5:4 In the name of our Lord Jesus Christ, when ye are sens eum qui sic operatus est 5:4 in nomine Domini nostri

gathered together, and my spirit, with the power of our Lord Jesus Christ,

Iesu Christi congregatis uobis et meo spiritu cum uirtute Do-

5:5 To deliver such an one unto Satan for the destruction mini Iesu 5:5 tradere huiusmodi Satanae in interitum carnis

of the flesh, that the spirit may be saved in the day of the Lord Jesus.

ut spiritus saluus sit in die Domini Iesu 5:6 non bona glo-

5:6 Your glorying is not good. Know ye not that a little leaven leaveneth the whole lump?

riatio uestra nescitis quia modicum fermentum totam mas-

5:7 Purge out therefore the old leaven, that ye may be sam corrumpit 5:7 expurgate uetus fermentum ut sitis noua

a new lump, as ye are unleavened. For even Christ our passover is sacrificed for us:

consparsio sicut estis azymi etenim pascha nostrum immola-

5:8 Therefore let us keep the feast, not with old leaven, tus est Christus 5:8 itaque epulemur non in fermento ueteri

neither with the leaven of malice and wickedness; but with the unleavened bread of sincerity and truth.

neque in fermento malitiae et nequitiae sed in azymis since-

5:9 I wrote unto you in an epistle not to company with ritatis et ueritatis 5:9 scripsi uobis in epistula ne commis-

fornicators:

5:10 Yet not altogether with the fornicators of this world, ceamini fornicariis 5:10 non utique fornicariis huius mundi

or with the covetous, or extortioners, or with idolaters; for then must ye needs go out of the world.

aut auaris aut rapacibus aut idolis seruientibus alioquin de-

5:11 But now I have written unto you not to keep com- bueratis de hoc mundo exisse 5:11 nunc autem scripsi uobis

pany, if any man that is called a brother be a fornicator, or covetous, or an idolater, or a railer, or a drunkard, or an

extortioner; with such an one no not to eat. non commisceri si is qui frater nominatur est fornicator aut

auarus aut idolis seruiens aut maledicus aut ebriosus aut ra-

5:12 For what have I to do to judge them also that are pax cum eiusmodi nec cibum sumere 5:12 quid enim mihi

without? do not ye judge them that are within?

de his qui foris sunt iudicare nonne de his qui intus sunt uos

5:13 But them that are without God judgeth. Therefore put iudicatis 5:13 nam eos qui foris sunt Deus iudicabit auferte

away from among yourselves that wicked person.

malum ex uobis ipsis

6:1 Dare any of you, having a matter against another, go 6:1 audet aliquis uestrum habens negotium aduersus alte-

to law before the unjust, and not before the saints?

6:2 Do ye not know that the saints shall judge the world? rum iudicari apud iniquos et non apud sanctos 6:2 an nesci-

and if the world shall be judged by you, are ye unworthy to judge the smallest matters?

tis quoniam sancti de mundo iudicabunt et si in uobis iudica-

6:3 Know ye not that we shall judge angels? how much bitur mundus indigni estis qui de minimis iudicetis 6:3 ne-

more things that pertain to this life?

scitis quoniam angelos iudicabimus quanto magis saecularia

6:4 If then ye have judgments of things pertaining to 6:4 saecularia igitur iudicia si habueritis contemptibiles qui

this life, set them to judge who are least esteemed in the church.

sunt in ecclesia illos constituite ad iudicandum 6:5 ad uere-

6:5 I speak to your shame. Is it so, that there is not a wise man among you? no, not one that shall be able to judge between his brethren?

cundiam uestram dico sic non est inter uos sapiens quisquam

6:6 But brother goeth to law with brother, and that before qui possit iudicare inter fratrem suum 6:6 sed frater cum the unbelievers.

6:7 Now therefore there is utterly a fault among you, be- fratre iudicio contendit et hoc apud infideles 6:7 iam qui-

cause ye go to law one with another. Why do ye not rather take wrong? why do ye not rather suffer yourselves to be

defrauded? dem omnino delictum est in uobis quod iudicia habetis in-

ter uos quare non magis iniuriam accipitis quare non magis

6:8 Nay, ye do wrong, and defraud, and that your brethren. fraudem patimini 6:8 sed uos iniuriam facitis et fraudatis

Epistola Pauli Ad Corinthios I

et hoc fratribus 6:9 an nescitis quia iniqui regnum Dei non 6:9 Know ye not that the unrighteous shall not inherit the

kingdom of God? Be not deceived: neither fornicators, nor idolaters, nor adulterers, nor effeminate, nor abusers

possidebunt nolite errare neque fornicarii neque idolis ser- of themselves with mankind,

uientes neque adulteri 6:10 neque molles neque masculo- 6:10 Nor thieves, nor covetous, nor drunkards, nor revil-

ers, nor extortioners, shall inherit the kingdom of God.

rum concubitores neque fures neque auari neque ebriosi neque maledici neque rapaces regnum Dei possidebunt 6:11 6:11 And such were some of you: but ye are washed, but

ye are sanctified, but ye are justified in the name of the Lord Jesus, and by the Spirit of our God.

et haec quidam fuistis sed abluti estis sed sanctificati estis sed iustificati estis in nomine Domini nostri Iesu Christi et in Spiritu Dei nostri 6:12 omnia mihi licent sed non omnia 6:12 All things are lawful unto me, but all things are not

expedient: all things are lawful for me, but I will not be brought under the power of any.

expediunt omnia mihi licent sed ego sub nullius rediiar potestate 6:13 esca uentri et uenter escis Deus autem et hunc et 6:13 Meats for the belly, and the belly for meats: but God

shall destroy both it and them. Now the body is not for fornication, but for the Lord; and the Lord for the body.

haec destruet corpus autem non fornicationi sed Domino et Dominus corpori 6:14 Deus uero et Dominum suscitauit et 6:14 And God hath both raised up the Lord, and will also

raise up us by his own power.

nos suscitabit per uirtutem suam 6:15 nescitis quoniam cor- 6:15 Know ye not that your bodies are the members of

Christ? shall I then take the members of Christ, and make them the members of an harlot? God forbid.

pora uestra membra Christi sunt tollens ergo membra Christi faciam membra meretricis absit 6:16 an nescitis quoniam 6:16 What? know ye not that he which is joined to an

harlot is one body? for two, saith he, shall be one flesh.

qui adheret meretrici unum corpus efficitur erunt enim inquit duo in carne una 6:17 qui autem adheret Domino unus spi- 6:17 But he that is joined unto the Lord is one spirit. ritus est 6:18 fugite fornicationem omne peccatum quod- 6:18 Flee fornication. Every sin that a man doeth is with-

out the body; but he that committeth fornication sinneth against his own body.

cumque fecerit homo extra corpus est qui autem fornicatur in corpus suum peccat 6:19 an nescitis quoniam membra 6:19 What? know ye not that your body is the temple of

the Holy Ghost which is in you, which ye have of God, and ye are not your own?

uestra templum est Spiritus Sancti qui in uobis est quem habetis a Deo et non estis uestri 6:20 empti enim estis pretio 6:20 For ye are bought with a price: therefore glorify God

in your body, and in your spirit, which are God’s.

magno glorificate et portate Deum in corpore uestro

7:1 de quibus autem scripsistis bonum est homini mulie- 7:1 Now concerning the things whereof ye wrote unto me:

It is good for a man not to touch a woman.

rem non tangere 7:2 propter fornicationes autem unusquis- 7:2 Nevertheless, to avoid fornication, let every man have

his own wife, and let every woman have her own husband.

que suam uxorem habeat et unaquaeque suum uirum habeat 7:3 uxori uir debitum reddat similiter autem et uxor uiro 7:4 7:3 Let the husband render unto the wife due benevolence:

and likewise also the wife unto the husband.

7:4 The wife hath not power of her own body, but the husband: and likewise also the husband hath not power of his own body, but the wife.

mulier sui corporis potestatem non habet sed uir similiter autem et uir sui corporis potestatem non habet sed mulier 7:5 7:5 Defraud ye not one the other, except it be with con-

sent for a time, that ye may give yourselves to fasting and prayer; and come together again, that Satan tempt you not

nolite fraudare inuicem nisi forte ex consensu ad tempus ut for your incontinency. uacetis orationi et iterum reuertimini in id ipsum ne temptet uos Satanas propter incontinentiam uestram 7:6 hoc autem 7:6 But I speak this by permission, and not of command-

ment.

dico secundum indulgentiam non secundum imperium 7:7 7:7 For I would that all men were even as I myself. But ev-

ery man hath his proper gift of God, one after this manner, and another after that.

Biblia Sacra Vulgata

uolo autem omnes homines esse sicut me ipsum sed unusquisque proprium habet donum ex Deo alius quidem sic al-

7:8 I say therefore to the unmarried and widows, It is good ius uero sic 7:8 dico autem non nuptis et uiduis bonum est for them if they abide even as I.

7:9 But if they cannot contain, let them marry: for it is illis si sic maneant sicut et ego 7:9 quod si non se continent

better to marry than to burn.

7:10 And unto the married I command, yet not I, but the nubant melius est enim nubere quam uri 7:10 his autem qui

Lord, Let not the wife depart from her husband:

matrimonio iuncti sunt praecipio non ego sed Dominus uxo-

7:11 But and if she depart, let her remain unmarried or rem a uiro non discedere 7:11 quod si discesserit manere

be reconciled to her husband: and let not the husband put away his wife.

innuptam aut uiro suo reconciliari et uir uxorem ne dimittat

7:12 But to the rest speak I, not the Lord: If any brother 7:12 nam ceteris ego dico non Dominus si quis frater uxo-

hath a wife that believeth not, and she be pleased to dwell with him, let him not put her away.

rem habet infidelem et haec consentit habitare cum illo non

7:13 And the woman which hath an husband that believeth dimittat illam 7:13 et si qua mulier habet uirum infidelem

not, and if he be pleased to dwell with her, let her not leave him.

et hic consentit habitare cum illa non dimittat uirum 7:14

7:14 For the unbelieving husband is sanctified by the wife, and the unbelieving wife is sanctified by the husband: else were your children unclean; but now are they holy.

sanctificatus est enim uir infidelis in muliere fideli et sanctificata est mulier infidelis per uirum fidelem alioquin filii

7:15 But if the unbelieving depart, let him depart. A uestri inmundi essent nunc autem sancti sunt 7:15 quod si

brother or a sister is not under bondage in such cases: but God hath called us to peace.

infidelis discedit discedat non est enim seruituti subiectus frater aut soror in eiusmodi in pace autem uocauit nos De-

7:16 For what knowest thou, O wife, whether thou shalt us 7:16 unde enim scis mulier si uirum saluum facies aut

save thy husband? or how knowest thou, O man, whether thou shalt save thy wife?

unde scis uir si mulierem saluam facies 7:17 nisi unicuique

7:17 But as God hath distributed to every man, as the Lord hath called every one, so let him walk. And so ordain I in all churches.

sicut diuisit Dominus unumquemque sicut uocauit Deus ita

7:18 Is any man called being circumcised? let him not ambulet et sic in omnibus ecclesiis doceo 7:18 circumcisus

become uncircumcised. Is any called in uncircumcision? let him not be circumcised.

aliquis uocatus est non adducat praeputium in praeputio ali-

7:19 Circumcision is nothing, and uncircumcision is noth- quis uocatus est non circumcidatur 7:19 circumcisio nihil ing, but the keeping of the commandments of God.

est et praeputium nihil est sed obseruatio mandatorum Dei

7:20 Let every man abide in the same calling wherein he 7:20 unusquisque in qua uocatione uocatus est in ea perma-

was called. 7:21 Art thou called being a servant? care not for it: but if neat 7:21 seruus uocatus es non sit tibi curae sed et si pothou mayest be made free, use it rather.

7:22 For he that is called in the Lord, being a servant, is tes liber fieri magis utere 7:22 qui enim in Domino uocatus

the Lord’s freeman: likewise also he that is called, being free, is Christ’s servant.

est seruus libertus est Domini similiter qui liber uocatus est

7:23 Ye are bought with a price; be not ye the servants of seruus est Christi 7:23 pretio empti estis nolite fieri serui

men. 7:24 Brethren, let every man, wherein he is called, therein hominum 7:24 unusquisque in quo uocatus est fratres in abide with God.

7:25 Now concerning virgins I have no commandment of hoc maneat apud Deum 7:25 de uirginibus autem praece-

the Lord: yet I give my judgment, as one that hath obtained mercy of the Lord to be faithful.

ptum Domini non habeo consilium autem do tamquam mi-

7:26 I suppose therefore that this is good for the present sericordiam consecutus a Domino ut sim fidelis 7:26 exisdistress, I say, that it is good for a man so to be.

Epistola Pauli Ad Corinthios I

timo ergo hoc bonum esse propter instantem necessitatem quoniam bonum est homini sic esse 7:27 alligatus es ux- 7:27 Art thou bound unto a wife? seek not to be loosed.

Art thou loosed from a wife? seek not a wife.

ori noli quaerere solutionem solutus es ab uxore noli quaerere uxorem 7:28 si autem acceperis uxorem non peccasti 7:28 But and if thou marry, thou hast not sinned; and if a

virgin marry, she hath not sinned. Nevertheless such shall have trouble in the flesh: but I spare you.

et si nupserit uirgo non peccauit tribulationem tamen carnis

habebunt huiusmodi ego autem uobis parco 7:29 hoc ita- 7:29 But this I say, brethren, the time is short: it remaineth,

that both they that have wives be as though they had none;

que dico fratres tempus breue est reliquum est ut qui habent uxores tamquam non habentes sint 7:30 et qui flent tam- 7:30 And they that weep, as though they wept not; and

they that rejoice, as though they rejoiced not; and they that buy, as though they possessed not;

quam non flentes et qui gaudent tamquam non gaudentes et qui emunt tamquam non possidentes 7:31 et qui utuntur 7:31 And they that use this world, as not abusing it: for

the fashion of this world passeth away.

hoc mundo tamquam non utantur praeterit enim figura huius mundi 7:32 uolo autem uos sine sollicitudine esse qui sine 7:32 But I would have you without carefulness. He that

is unmarried careth for the things that belong to the Lord, how he may please the Lord:

uxore est sollicitus est quae Domini sunt quomodo placeat

Deo 7:33 qui autem cum uxore est sollicitus est quae sunt 7:33 But he that is married careth for the things that are of

the world, how he may please his wife.

mundi quomodo placeat uxori et diuisus est 7:34 et mulier 7:34 There is difference also between a wife and a virgin.

The unmarried woman careth for the things of the Lord, that she may be holy both in body and in spirit: but she that is married careth for the things of the world, how she may please her husband.

innupta et uirgo cogitat quae Domini sunt ut sit sancta et corpore et spiritu quae autem nupta est cogitat quae sunt mundi quomodo placeat uiro 7:35 porro hoc ad utilitatem 7:35 And this I speak for your own profit; not that I may

cast a snare upon you, but for that which is comely, and that ye may attend upon the Lord without distraction.

uestram dico non ut laqueum uobis iniciam sed ad id quod honestum est et quod facultatem praebeat sine inpedimento Dominum obseruandi 7:36 si quis autem turpem se uideri 7:36 But if any man think that he behaveth himself un-

comely toward his virgin, if she pass the flower of her age, and need so require, let him do what he will, he sinneth

existimat super uirgine sua quod sit superadulta et ita oportet not: let them marry. fieri quod uult faciat non peccat nubat 7:37 nam qui statuit 7:37 Nevertheless he that standeth stedfast in his heart,

having no necessity, but hath power over his own will, and hath so decreed in his heart that he will keep his virgin,

in corde suo firmus non habens necessitatem potestatem au- doeth well. tem habet suae uoluntatis et hoc iudicauit in corde suo seruare uirginem suam bene facit 7:38 igitur et qui matrimo- 7:38 So then he that giveth her in marriage doeth well; but

he that giveth her not in marriage doeth better.

nio iungit uirginem suam bene facit et qui non iungit melius facit 7:39 mulier alligata est quanto tempore uir eius uiuit 7:39 The wife is bound by the law as long as her husband

liveth; but if her husband be dead, she is at liberty to be married to whom she will; only in the Lord.

quod si dormierit uir eius liberata est cui uult nubat tantum in Domino 7:40 beatior autem erit si sic permanserit secun- 7:40 But she is happier if she so abide, after my judgment:

and I think also that I have the Spirit of God.

dum meum consilium puto autem quod et ego Spiritum Dei habeo

8:1 de his autem quae idolis sacrificantur scimus quia om- 8:1 Now as touching things offered unto idols, we know

that we all have knowledge. Knowledge puffeth up, but charity edifieth.

Biblia Sacra Vulgata

nes scientiam habemus scientia inflat caritas uero aedificat

8:2 And if any man think that he knoweth any thing, he 8:2 si quis se existimat scire aliquid nondum cognouit que-

knoweth nothing yet as he ought to know.

8:3 But if any man love God, the same is known of him. madmodum oporteat eum scire 8:3 si quis autem diligit De-

8:4 As concerning therefore the eating of those things that um hic cognitus est ab eo 8:4 de escis autem quae idolis are offered in sacrifice unto idols, we know that an idol is nothing in the world, and that there is none other God but

one. immolantur scimus quia nihil est idolum in mundo et quod

8:5 For though there be that are called gods, whether in nullus Deus nisi unus 8:5 nam et si sunt qui dicantur dii heaven or in earth, (as there be gods many, and lords many,)

siue in caelo siue in terra siquidem sunt dii multi et domini

8:6 But to us there is but one God, the Father, of whom are multi 8:6 nobis tamen unus Deus Pater ex quo omnia et nos all things, and we in him; and one Lord Jesus Christ, by whom are all things, and we by him.

in illum et unus Dominus Iesus Christus per quem omnia et

8:7 Howbeit there is not in every man that knowledge: for nos per ipsum 8:7 sed non in omnibus est scientia quidam some with conscience of the idol unto this hour eat it as a thing offered unto an idol; and their conscience being

weak is defiled. autem conscientia usque nunc idoli quasi idolothytum man-8:8 But meat commendeth us not to God: for neither, if ducant et conscientia ipsorum cum sit infirma polluitur 8:8

we eat, are we the better; neither, if we eat not, are we the worse.

esca autem nos non commendat Deo neque si non manducauerimus deficiemus neque si manducauerimus abundabi-

8:9 But take heed lest by any means this liberty of your’s mus 8:9 uidete autem ne forte haec licentia uestra offen-

become a stumblingblock to them that are weak.

8:10 For if any man see thee which hast knowledge sit at diculum fiat infirmibus 8:10 si enim quis uiderit eum qui

meat in the idol’s temple, shall not the conscience of him which is weak be emboldened to eat those things which

are offered to idols; habet scientiam in idolio recumbentem nonne conscientia

eius cum sit infirma aedificabitur ad manducandum idolo-

8:11 And through thy knowledge shall the weak brother thyta 8:11 et peribit infirmus in tua scientia frater propter

perish, for whom Christ died?

8:12 But when ye sin so against the brethren, and wound quem Christus mortuus est 8:12 sic autem peccantes in fra-

their weak conscience, ye sin against Christ.

tres et percutientes conscientiam eorum infirmam in Christo

8:13 Wherefore, if meat make my brother to offend, I peccatis 8:13 quapropter si esca scandalizat fratrem meum

will eat no flesh while the world standeth, lest I make my brother to offend.

non manducabo carnem in aeternum ne fratrem meum scandalizem

9:1 Am I not an apostle? am I not free? have I not seen 9:1 non sum liber non sum apostolus nonne Iesum Domi-

Jesus Christ our Lord? are not ye my work in the Lord? 9:2 If I be not an apostle unto others, yet doubtless I am to num nostrum uidi non opus meum uos estis in Domino 9:2 you: for the seal of mine apostleship are ye in the Lord.

si aliis non sum apostolus sed tamen uobis sum nam signacu-

9:3 Mine answer to them that do examine me is this, lum apostolatus mei uos estis in Domino 9:3 mea defensio 9:4 Have we not power to eat and to drink? apud eos qui me interrogant haec est 9:4 numquid non ha-9:5 Have we not power to lead about a sister, a wife, as bemus potestatem manducandi et bibendi 9:5 numquid non

well as other apostles, and as the brethren of the Lord, and Cephas?

habemus potestatem sororem mulierem circumducendi sicut

9:6 Or I only and Barnabas, have not we power to forbear et ceteri apostoli et fratres Domini et Cephas 9:6 aut solus

working?

9:7 Who goeth a warfare any time at his own charges? who ego et Barnabas non habemus potestatem hoc operandi 9:7

planteth a vineyard, and eateth not of the fruit thereof? or who feedeth a flock, and eateth not of the milk of the flock?

Epistola Pauli Ad Corinthios I

quis militat suis stipendiis umquam quis plantat uineam et fructum eius non edit quis pascit gregem et de lacte gregis non manducat 9:8 numquid secundum hominem haec dico 9:8 Say I these things as a man? or saith not the law the

same also?

an et lex haec non dicit 9:9 scriptum est enim in lege Mosi 9:9 For it is written in the law of Moses, Thou shalt not

muzzle the mouth of the ox that treadeth out the corn. Doth God take care for oxen?

non alligabis os boui trituranti numquid de bubus cura est

Deo 9:10 an propter nos utique dicit nam propter nos sc- 9:10 Or saith he it altogether for our sakes? For our sakes, no doubt, this is written: that he that ploweth should plow in hope; and that he that thresheth in hope should be par-

ripta sunt quoniam debet in spe qui arat arare et qui triturat taker of his hope. in spe fructus percipiendi 9:11 si nos uobis spiritalia semi- 9:11 If we have sown unto you spiritual things, is it a great

thing if we shall reap your carnal things?

nauimus magnum est si nos carnalia uestra metamus 9:12 si 9:12 If others be partakers of this power over you, are not

we rather? Nevertheless we have not used this power; but suffer all things, lest we should hinder the gospel of Christ.

alii potestatis uestrae participes sunt non potius nos sed non usi sumus hac potestate sed omnia sustinemus ne quod offendiculum demus euangelio Christi 9:13 nescitis quoniam 9:13 Do ye not know that they which minister about holy

things live of the things of the temple? and they which wait at the altar are partakers with the altar?

qui in sacrario operantur quae de sacrario sunt edunt qui altario deseruiunt cum altario participantur 9:14 ita et Dominus 9:14 Even so hath the Lord ordained that they which

preach the gospel should live of the gospel.

ordinauit his qui euangelium adnuntiant de euangelio uiuere 9:15 ego autem nullo horum usus sum non scripsi autem haec 9:15 But I have used none of these things: neither have I

written these things, that it should be so done unto me: for it were better for me to die, than that any man should make

ut ita fiant in me bonum est enim mihi magis mori quam my glorying void. ut gloriam meam quis euacuet 9:16 nam si euangelizauero 9:16 For though I preach the gospel, I have nothing to

glory of: for necessity is laid upon me; yea, woe is unto me, if I preach not the gospel!

non est mihi gloria necessitas enim mihi incumbit uae enim

mihi est si non euangelizauero 9:17 si enim uolens hoc ago 9:17 For if I do this thing willingly, I have a reward: but if against my will, a dispensation of the gospel is committed unto me.

mercedem habeo si autem inuitus dispensatio mihi credita est 9:18 quae est ergo merces mea ut euangelium praedi- 9:18 What is my reward then? Verily that, when I preach

the gospel, I may make the gospel of Christ without charge, that I abuse not my power in the gospel.

cans sine sumptu ponam euangelium ut non abutar potestate mea in euangelio 9:19 nam cum liber essem ex omnibus 9:19 For though I be free from all men, yet have I made

myself servant unto all, that I might gain the more.

omnium me seruum feci ut plures lucri facerem 9:20 et fac- 9:20 And unto the Jews I became as a Jew, that I might

gain the Jews; to them that are under the law, as under the law, that I might gain them that are under the law;

tus sum Iudaeis tamquam Iudaeus ut Iudaeos lucrarer 9:21

9:21 To them that are without law, as without law, (being not without law to God, but under the law to Christ,) that I might gain them that are without law.

his qui sub lege sunt quasi sub lege essem cum ipse non essem sub lege ut eos qui sub lege erant lucri facerem his qui sine lege erant tamquam sine lege essem cum sine lege Dei non essem sed in lege essem Christi ut lucri facerem eos qui sine lege erant 9:22 factus sum infirmis infirmus ut infirmos 9:22 To the weak became I as weak, that I might gain the

weak: I am made all things to all men, that I might by all means save some.

lucri facerem omnibus omnia factus sum ut omnes facerem saluos 9:23 omnia autem facio propter euangelium ut parti- 9:23 And this I do for the gospel’s sake, that I might be

partaker thereof with you.

Biblia Sacra Vulgata

9:24 Know ye not that they which run in a race run all, but ceps eius efficiar 9:24 nescitis quod hii qui in stadio currunt

one receiveth the prize? So run, that ye may obtain.

omnes quidem currunt sed unus accipit brauium sic currite ut

9:25 And every man that striveth for the mastery is tem- conprehendatis 9:25 omnis autem qui in agone contendit ab

perate in all things. Now they do it to obtain a corruptible crown; but we an incorruptible.

omnibus se abstinet et illi quidem ut corruptibilem coronam

9:26 I therefore so run, not as uncertainly; so fight I, not accipiant nos autem incorruptam 9:26 ego igitur sic curro

as one that beateth the air:

non quasi in incertum sic pugno non quasi aerem uerberans

9:27 But I keep under my body, and bring it into subjec- 9:27 sed castigo corpus meum et in seruitutem rediio ne forte

tion: lest that by any means, when I have preached to others, I myself should be a castaway.

cum aliis praedicauerim ipse reprobus efficiar

10:1 Moreover, brethren, I would not that ye should be 10:1 nolo enim uos ignorare fratres quoniam patres nos-

ignorant, how that all our fathers were under the cloud, and all passed through the sea;

tri omnes sub nube fuerunt et omnes mare transierunt 10:2

10:2 And were all baptized unto Moses in the cloud and in the sea;

10:3 And did all eat the same spiritual meat; et omnes in Mose baptizati sunt in nube et in mari 10:3

10:4 And did all drink the same spiritual drink: for they et omnes eandem escam spiritalem manducauerunt 10:4 et drank of that spiritual Rock that followed them: and that Rock was Christ.

omnes eundem potum spiritalem biberunt bibebant autem de spiritali consequenti eos petra petra autem erat Christus

10:5 But with many of them God was not well pleased: for 10:5 sed non in pluribus eorum beneplacitum est Deo nam

they were overthrown in the wilderness. 10:6 Now these things were our examples, to the intent we prostrati sunt in deserto 10:6 haec autem in figura facta should not lust after evil things, as they also lusted.

sunt nostri ut non simus concupiscentes malorum sicut et

10:7 Neither be ye idolaters, as were some of them; as it illi concupierunt 10:7 neque idolorum cultores efficiamini

is written, The people sat down to eat and drink, and rose up to play.

sicut quidam ex ipsis quemadmodum scriptum est sedit po-

10:8 Neither let us commit fornication, as some of them pulus manducare et bibere et surrexerunt ludere 10:8 neque

committed, and fell in one day three and twenty thousand.

fornicemur sicut quidam ex ipsis fornicati sunt et ceciderunt

10:9 Neither let us tempt Christ, as some of them also una die uiginti tria milia 10:9 neque temptemus Christum

tempted, and were destroyed of serpents.

sicut quidam eorum temptauerunt et a serpentibus perierunt

10:10 Neither murmur ye, as some of them also mur- 10:10 neque murmuraueritis sicut quidam eorum murmura-

mured, and were destroyed of the destroyer.

10:11 Now all these things happened unto them for ensam- uerunt et perierunt ab exterminatore 10:11 haec autem om-

ples: and they are written for our admonition, upon whom the ends of the world are come.

nia in figura contingebant illis scripta sunt autem ad correp-

10:12 Wherefore let him that thinketh he standeth take tionem nostram in quos fines saeculorum deuenerunt 10:12

heed lest he fall.

10:13 There hath no temptation taken you but such as is itaque qui se existimat stare uideat ne cadat 10:13 temp-

common to man: but God is faithful, who will not suffer you to be tempted above that ye are able; but will with the temptation also make a way to escape, that ye may be able to bear it.

tatio uos non adprehendat nisi humana fidelis autem Deus qui non patietur uos temptari super id quod potestis sed faciet cum temptatione etiam prouentum ut possitis sustinere

10:14 Wherefore, my dearly beloved, flee from idolatry. 10:14 propter quod carissimi mihi fugite ab idolorum cultura 10:15 I speak as to wise men; judge ye what I say. 10:15 ut prudentibus loquor uos iudicate quod dico 10:16 ca-

10:16 The cup of blessing which we bless, is it not the communion of the blood of Christ? The bread which we break, is it not the communion of the body of Christ?

Epistola Pauli Ad Corinthios I

licem benedictionis cui benedicimus nonne communicatio sanguinis Christi est et panis quem frangimus nonne participatio corporis Domini est 10:17 quoniam unus panis unum 10:17 For we being many are one bread, and one body: for

we are all partakers of that one bread.

corpus multi sumus omnes quidem de uno pane participamur 10:18 uidete Israhel secundum carnem nonne qui edunt hos- 10:18 Behold Israel after the flesh: are not they which eat

of the sacrifices partakers of the altar?

tias participes sunt altaris 10:19 quid ergo dico quod ido- 10:19 What say I then? that the idol is any thing, or that

which is offered in sacrifice to idols is any thing?

lis immolatum sit aliquid aut quod idolum sit aliquid 10:20 10:20 But I say, that the things which the Gentiles sacri-

fice, they sacrifice to devils, and not to God: and I would not that ye should have fellowship with devils.

sed quae immolant gentes daemoniis immolant et non Deo nolo autem uos socios fieri daemoniorum non potestis calicem Domini bibere et calicem daemoniorum 10:21 non po- 10:21 Ye cannot drink the cup of the Lord, and the cup of

devils: ye cannot be partakers of the Lord’s table, and of the table of devils.

testis mensae Domini participes esse et mensae daemoniorum 10:22 an aemulamur Dominum numquid fortiores illo 10:22 Do we provoke the Lord to jealousy? are we

stronger than he?

sumus omnia licent sed non omnia expediunt 10:23 omnia 10:23 All things are lawful for me, but all things are not expedient: all things are lawful for me, but all things edify not.

licent sed non omnia aedificant 10:24 nemo quod suum est

10:24 Let no man seek his own, but every man another’s wealth.

quaerat sed quod alterius 10:25 omne quod in macello uenit 10:25 Whatsoever is sold in the shambles, that eat, asking

no question for conscience sake:

manducate nihil interrogantes propter conscientiam 10:26 10:26 For the earth is the Lord’s, and the fulness thereof. Domini est terra et plenitudo eius 10:27 si quis uocat uos 10:27 If any of them that believe not bid you to a feast,

and ye be disposed to go; whatsoever is set before you, eat, asking no question for conscience sake.

infidelium et uultis ire omne quod uobis adponitur manducate nihil interrogantes propter conscientiam 10:28 si quis 10:28 But if any man say unto you, This is offered in sac-

rifice unto idols, eat not for his sake that shewed it, and for conscience sake: for the earth is the Lord’s, and the

autem dixerit hoc immolaticium est idolis nolite manducare fulness thereof: propter illum qui indicauit et propter conscientiam 10:29 10:29 Conscience, I say, not thine own, but of the other:

for why is my liberty judged of another man’s conscience?

conscientiam autem dico non tuam sed alterius ut quid enim libertas mea iudicatur ab alia conscientia 10:30 si ego cum 10:30 For if I by grace be a partaker, why am I evil spoken

of for that for which I give thanks?

gratia participo quid blasphemor pro eo quod gratias ago 10:31 siue ergo manducatis siue bibitis uel aliud quid faci- 10:31 Whether therefore ye eat, or drink, or whatsoever ye

do, do all to the glory of God.

tis omnia in gloriam Dei facite 10:32 sine offensione estote 10:32 Give none offence, neither to the Jews, nor to the

Gentiles, nor to the church of God:

Iudaeis et gentilibus et ecclesiae Dei 10:33 sicut et ego per 10:33 Even as I please all men in all things, not seeking

mine own profit, but the profit of many, that they may be saved.

omnia omnibus placeo non quaerens quod mihi utile est sed quod multis ut salui fiant

11:1 imitatores mei estote sicut et ego Christi 11:2 laudo 11:1 Be ye followers of me, even as I also am of Christ.

11:2 Now I praise you, brethren, that ye remember me in all things, and keep the ordinances, as I delivered them to you.

autem uos fratres quod omnia mei memores estis et sicut tradidi uobis praecepta mea tenetis 11:3 uolo autem uos scire 11:3 But I would have you know, that the head of every

man is Christ; and the head of the woman is the man; and the head of Christ is God.

quod omnis uiri caput Christus est caput autem mulieris uir 11:4 Every man praying or prophesying, having his head caput uero Christi Deus 11:4 omnis uir orans aut prophe-

covered, dishonoureth his head. 11:5 But every woman that prayeth or prophesieth with her tans uelato capite deturpat caput suum 11:5 omnis autem head uncovered dishonoureth her head: for that is even all one as if she were shaven.

mulier orans aut prophetans non uelato capite deturpat caput

11:6 For if the woman be not covered, let her also be suum unum est enim atque si decaluetur 11:6 nam si non

shorn: but if it be a shame for a woman to be shorn or shaven, let her be covered.

uelatur mulier et tondeatur si uero turpe est mulieri tonderi

11:7 For a man indeed ought not to cover his head, foras- aut decaluari uelet caput suum 11:7 uir quidem non debet

much as he is the image and glory of God: but the woman is the glory of the man.

uelare caput quoniam imago et gloria est Dei mulier autem

11:8 For the man is not of the woman: but the woman of gloria uiri est 11:8 non enim uir ex muliere est sed mulier

the man.

11:9 Neither was the man created for the woman; but the ex uiro 11:9 etenim non est creatus uir propter mulierem

woman for the man.

11:10 For this cause ought the woman to have power on sed mulier propter uirum 11:10 ideo debet mulier potesta-

her head because of the angels. 11:11 Nevertheless neither is the man without the woman, tem habere supra caput propter angelos 11:11 uerumtamen neither the woman without the man, in the Lord.

neque uir sine muliere neque mulier sine uiro in Domino

11:12 For as the woman is of the man, even so is the man 11:12 nam sicut mulier de uiro ita et uir per mulierem om-

also by the woman; but all things of God. 11:13 Judge in yourselves: is it comely that a woman pray nia autem ex Deo 11:13 uos ipsi iudicate decet mulierem unto God uncovered? 11:14 Doth not even nature itself teach you, that, if a man non uelatam orare Deum 11:14 nec ipsa natura docet uos have long hair, it is a shame unto him?

11:15 But if a woman have long hair, it is a glory to her: quod uir quidem si comam nutriat ignominia est illi 11:15

for her hair is given her for a covering.

mulier uero si comam nutriat gloria est illi quoniam capilli

11:16 But if any man seem to be contentious, we have no pro uelamine ei dati sunt 11:16 si quis autem uidetur con-

such custom, neither the churches of God.

tentiosus esse nos talem consuetudinem non habemus neque

11:17 Now in this that I declare unto you I praise you not, ecclesiae Dei 11:17 hoc autem praecipio non laudans quod

that ye come together not for the better, but for the worse.

11:18 For first of all, when ye come together in the church, non in melius sed in deterius conuenitis 11:18 primum qui-

I hear that there be divisions among you; and I partly believe it.

dem conuenientibus uobis in ecclesia audio scissuras esse

11:19 For there must be also heresies among you, that they et ex parte credo 11:19 nam oportet et hereses esse ut et which are approved may be made manifest among you.

11:20 When ye come together therefore into one place, qui probati sunt manifesti fiant in uobis 11:20 conuenien-

this is not to eat the Lord’s supper.

tibus ergo uobis in unum iam non est dominicam cenam

11:21 For in eating every one taketh before other his own manducare 11:21 unusquisque enim suam cenam praesumit supper: and one is hungry, and another is drunken.

ad manducandum et alius quidem esurit alius autem ebrius

11:22 What? have ye not houses to eat and to drink in? or est 11:22 numquid domos non habetis ad manducandum et despise ye the church of God, and shame them that have not? What shall I say to you? shall I praise you in this? I

praise you not. bibendum aut ecclesiam Dei contemnitis et confunditis eos qui non habent quid dicam uobis laudo uos in hoc non laudo

11:23 For I have received of the Lord that which also I 11:23 ego enim accepi a Domino quod et tradidi uobis quo-

delivered unto you, That the Lord Jesus the same night in which he was betrayed took bread:

niam Dominus Iesus in qua nocte tradebatur accepit panem

11:24 And when he had given thanks, he brake it, and said, 11:24 et gratias agens fregit et dixit hoc est corpus meum

Take, eat: this is my body, which is broken for you: this do in remembrance of me.

Epistola Pauli Ad Corinthios I

pro uobis hoc facite in meam commemorationem 11:25 si- 11:25 After the same manner also he took the cup, when

he had supped, saying, This cup is the new testament in my blood: this do ye, as oft as ye drink it, in remembrance

militer et calicem postquam cenauit dicens hic calix nouum of me. testamentum est in meo sanguine hoc facite quotienscumque

bibetis in meam commemorationem 11:26 quotienscumque 11:26 For as often as ye eat this bread, and drink this cup,

ye do shew the Lord’s death till he come.

enim manducabitis panem hunc et calicem bibetis mortem

Domini adnuntiatis donec ueniat 11:27 itaque quicumque 11:27 Wherefore whosoever shall eat this bread, and drink this cup of the Lord, unworthily, shall be guilty of the body and blood of the Lord.

manducauerit panem uel biberit calicem Domini indigne reus erit corporis et sanguinis Domini 11:28 probet autem 11:28 But let a man examine himself, and so let him eat of

that bread, and drink of that cup.

se ipsum homo et sic de pane illo edat et de calice bibat 11:29 qui enim manducat et bibit indigne iudicium sibi man- 11:29 For he that eateth and drinketh unworthily, eateth

and drinketh damnation to himself, not discerning the Lord’s body.

ducat et bibit non diiudicans corpus 11:30 ideo inter uos

11:30 For this cause many are weak and sickly among you, and many sleep.

multi infirmes et inbecilles et dormiunt multi 11:31 quod si _{11:31 For if we would judge ourselves, we should not be}

judged.

nosmet ipsos diiudicaremus non utique iudicaremur 11:32 11:32 But when we are judged, we are chastened of the

Lord, that we should not be condemned with the world.

dum iudicamur autem a Domino corripimur ut non cum hoc

mundo damnemur 11:33 itaque fratres mei cum conueni- 11:33 Wherefore, my brethren, when ye come together to

eat, tarry one for another.

tis ad manducandum inuicem expectate 11:34 si quis esurit 11:34 And if any man hunger, let him eat at home; that ye

come not together unto condemnation. And the rest will I set in order when I come.

domi manducet ut non in iudicium conueniatis cetera autem cum uenero disponam

12:1 de spiritalibus autem nolo uos ignorare fratres 12:2 12:1 Now concerning spiritual gifts, brethren, I would not have you ignorant. 12:2 Ye know that ye were Gentiles, carried away unto these dumb idols, even as ye were led.

scitis quoniam cum gentes essetis ad simulacra muta prout ducebamini euntes 12:3 ideo notum uobis facio quod nemo 12:3 Wherefore I give you to understand, that no man

speaking by the Spirit of God calleth Jesus accursed: and that no man can say that Jesus is the Lord, but by the Holy

in Spiritu Dei loquens dicit anathema Iesu et nemo potest Ghost. dicere Dominus Iesus nisi in Spiritu Sancto 12:4 diuisiones 12:4 Now there are diversities of gifts, but the same Spirit. uero gratiarum sunt idem autem Spiritus 12:5 et diuisiones 12:5 And there are differences of administrations, but the

same Lord.

ministrationum sunt idem autem Dominus 12:6 et diuisio- 12:6 And there are diversities of operations, but it is the

same God which worketh all in all.

nes operationum sunt idem uero Deus qui operatur omnia in omnibus 12:7 unicuique autem datur manifestatio Spiri- 12:7 But the manifestation of the Spirit is given to every

man to profit withal.

tus ad utilitatem 12:8 alii quidem per Spiritum datur sermo 12:8 For to one is given by the Spirit the word of wisdom;

to another the word of knowledge by the same Spirit;

sapientiae alii autem sermo scientiae secundum eundem Spiritum 12:9 alteri fides in eodem Spiritu alii gratia sanitatum 12:9 To another faith by the same Spirit; to another the

gifts of healing by the same Spirit;

in uno Spiritu 12:10 alii operatio uirtutum alii prophetatio 12:10 To another the working of miracles; to another

prophecy; to another discerning of spirits; to another divers kinds of tongues; to another the interpretation of

alii discretio spirituum alii genera linguarum alii interpreta- tongues: tio sermonum 12:11 haec autem omnia operatur unus atque 12:11 But all these worketh that one and the selfsame

Spirit, dividing to every man severally as he will.

Biblia Sacra Vulgata

12:12 For as the body is one, and hath many members, idem Spiritus diuidens singulis prout uult 12:12 sicut enim

and all the members of that one body, being many, are one body: so also is Christ.

corpus unum est et membra habet multa omnia autem membra corporis cum sint multa unum corpus sunt ita et Christus

12:13 For by one Spirit are we all baptized into one body, 12:13 etenim in uno Spiritu omnes nos in unum corpus bawhether we be Jews or Gentiles, whether we be bond or free; and have been all made to drink into one Spirit.

ptizati sumus siue Iudaei siue gentiles siue serui siue liberi

12:14 For the body is not one member, but many. et omnes unum Spiritum potati sumus 12:14 nam et cor-

12:15 If the foot shall say, Because I am not the hand, I pus non est unum membrum sed multa 12:15 si dixerit pes am not of the body; is it therefore not of the body?

quoniam non sum manus non sum de corpore non ideo non

12:16 And if the ear shall say, Because I am not the eye, I est de corpore 12:16 et si dixerit auris quia non sum oculus

am not of the body; is it therefore not of the body? 12:17 If the whole body were an eye, where were the hear- non sum de corpore non ideo non est de corpore 12:17 si ing? If the whole were hearing, where were the smelling?

totum corpus oculus ubi auditus si totum auditus ubi odo-

12:18 But now hath God set the members every one of ratus 12:18 nunc autem posuit Deus membra unumquod-

them in the body, as it hath pleased him.

12:19 And if they were all one member, where were the que eorum in corpore sicut uoluit 12:19 quod si essent om-

body? 12:20 But now are they many members, yet but one body. nia unum membrum ubi corpus 12:20 nunc autem multa

12:21 And the eye cannot say unto the hand, I have no quidem membra unum autem corpus 12:21 non potest di-

need of thee: nor again the head to the feet, I have no need of you.

cere oculus manui opera tua non indigeo aut iterum caput

12:22 Nay, much more those members of the body, which pedibus non estis mihi necessarii 12:22 sed multo magis

seem to be more feeble, are necessary:

quae uidentur membra corporis infirmiora esse necessariora

12:23 And those members of the body, which we think sunt 12:23 et quae putamus ignobiliora membra esse corpo-

to be less honourable, upon these we bestow more abundant honour; and our uncomely parts have more abundant

comeliness. ris his honorem abundantiorem circumdamus et quae inho-12:24 For our comely parts have no need: but God hath nesta sunt nostra abundantiorem honestatem habent 12:24

tempered the body together, having given more abundant honour to that part which lacked.

honesta autem nostra nullius egent sed Deus temperauit cor-

12:25 That there should be no schism in the body; but that pus ei cui deerat abundantiorem tribuendo honorem 12:25

the members should have the same care one for another.

ut non sit scisma in corpore sed id ipsum pro inuicem sol-

12:26 And whether one member suffer, all the members licita sint membra 12:26 et si quid patitur unum membrum

suffer with it; or one member be honoured, all the members rejoice with it.

conpatiuntur omnia membra siue gloriatur unum membrum

12:27 Now ye are the body of Christ, and members in par- congaudent omnia membra 12:27 uos autem estis corpus ticular. 12:28 And God hath set some in the church, first apostles, Christi et membra de membro 12:28 et quosdam quidem secondarily prophets, thirdly teachers, after that miracles, then gifts of healings, helps, governments, diversities of

tongues. posuit Deus in ecclesia primum apostolos secundo prophetas

tertio doctores deinde uirtutes exin gratias curationum opi-

12:29 Are all apostles? are all prophets? are all teachers? tulationes gubernationes genera linguarum 12:29 numquid are all workers of miracles?

omnes apostoli numquid omnes prophetae numquid omnes

12:30 Have all the gifts of healing? do all speak with doctores 12:30 numquid omnes uirtutes numquid omnes

tongues? do all interpret?

Epistola Pauli Ad Corinthios I

gratiam habent curationum numquid omnes linguis loquuntur numquid omnes interpretantur 12:31 aemulamini autem 12:31 But covet earnestly the best gifts: and yet shew I

unto you a more excellent way.

charismata maiora et adhuc excellentiorem uiam uobis demonstro

13:1 si linguis hominum loquar et angelorum caritatem au- 13:1 Though I speak with the tongues of men and of an-

gels, and have not charity, I am become as sounding brass, or a tinkling cymbal.

tem non habeam factus sum uelut aes sonans aut cymbalum tinniens 13:2 et si habuero prophetiam et nouerim my- 13:2 And though I have the gift of prophecy, and under-

stand all mysteries, and all knowledge; and though I have all faith, so that I could remove mountains, and have not

steria omnia et omnem scientiam et habuero omnem fidem charity, I am nothing. ita ut montes transferam caritatem autem non habuero nihil sum 13:3 et si distribuero in cibos pauperum omnes faculta- 13:3 And though I bestow all my goods to feed the poor,

and though I give my body to be burned, and have not charity, it profiteth me nothing.

tes meas et si tradidero corpus meum ut ardeam caritatem autem non habuero nihil mihi prodeet 13:4 caritas patiens est 13:4 Charity suffereth long, and is kind; charity envieth

not; charity vaunteth not itself, is not puffed up,

benigna est caritas non aemulatur non agit perperam non in-

flatur 13:5 non est ambitiosa non quaerit quae sua sunt non 13:5 Doth not behave itself unseemly, seeketh not her own,

is not easily provoked, thinketh no evil;

inritatur non cogitat malum 13:6 non gaudet super iniqui- 13:6 Rejoiceth not in iniquity, but rejoiceth in the truth; tatem congaudet autem ueritati 13:7 omnia suffert omnia 13:7 Beareth all things, believeth all things, hopeth all

things, endureth all things.

credit omnia sperat omnia sustinet 13:8 caritas numquam 13:8 Charity never faileth: but whether there be prophe-

cies, they shall fail; whether there be tongues, they shall cease; whether there be knowledge, it shall vanish away.

excidit siue prophetiae euacuabuntur siue linguae cessabunt siue scientia destruetur 13:9 ex parte enim cognoscimus et 13:9 For we know in part, and we prophesy in part. ex parte prophetamus 13:10 cum autem uenerit quod per- 13:10 But when that which is perfect is come, then that

which is in part shall be done away.

fectum est euacuabitur quod ex parte est 13:11 cum essem 13:11 When I was a child, I spake as a child, I understood

as a child, I thought as a child: but when I became a man, I put away childish things.

paruulus loquebar ut paruulus sapiebam ut paruulus cogitabam ut paruulus quando factus sum uir euacuaui quae erant paruuli 13:12 uidemus nunc per speculum in enigmate tunc 13:12 For now we see through a glass, darkly; but then

face to face: now I know in part; but then shall I know even as also I am known.

autem facie ad faciem nunc cognosco ex parte tunc autem

cognoscam sicut et cognitus sum 13:13 nunc autem manet 13:13 And now abideth faith, hope, charity, these three;

but the greatest of these is charity.

fides spes caritas tria haec maior autem his est caritas

14:1 sectamini caritatem aemulamini spiritalia magis au- 14:1 Follow after charity, and desire spiritual gifts, but

rather that ye may prophesy.

tem ut prophetetis 14:2 qui enim loquitur lingua non homi- 14:2 For he that speaketh in an unknown tongue speaketh

not unto men, but unto God: for no man understandeth him; howbeit in the spirit he speaketh mysteries.

nibus loquitur sed Deo nemo enim audit Spiritu autem loquitur mysteria 14:3 nam qui prophetat hominibus loquitur 14:3 But he that prophesieth speaketh unto men to edifica-

tion, and exhortation, and comfort.

aedificationem et exhortationem et consolationes 14:4 qui 14:4 He that speaketh in an unknown tongue edifieth him-

self; but he that prophesieth edifieth the church.

loquitur lingua semet ipsum aedificat qui autem prophetat

Biblia Sacra Vulgata

14:5 I would that ye all spake with tongues but rather that ecclesiam aedificat 14:5 uolo autem omnes uos loqui linye prophesied: for greater is he that prophesieth than he that speaketh with tongues, except he interpret, that the

church may receive edifying. guis magis autem prophetare nam maior est qui prophetat

quam qui loquitur linguis nisi si forte ut interpretetur ut ecc-

14:6 Now, brethren, if I come unto you speaking with lesia aedificationem accipiat 14:6 nunc autem fratres si ue-

tongues, what shall I profit you, except I shall speak to you either by revelation, or by knowledge, or by prophesying,

or by doctrine? nero ad uos linguis loquens quid uobis prodeeo nisi si uobis

loquar aut in reuelatione aut scientia aut prophetia aut in doc-

14:7 And even things without life giving sound, whether trina 14:7 tamen quae sine anima sunt uocem dantia siue pipe or harp, except they give a distinction in the sounds, how shall it be known what is piped or harped?

tibia siue cithara nisi distinctionem sonituum dederint quo-

14:8 For if the trumpet give an uncertain sound, who shall modo scietur quod canitur aut quod citharizatur 14:8 ete-

prepare himself to the battle?

nim si incertam uocem det tuba quis parabit se ad bellum

14:9 So likewise ye, except ye utter by the tongue words 14:9 ita et uos per linguam nisi manifestum sermonem dede-

easy to be understood, how shall it be known what is spoken? for ye shall speak into the air.

ritis quomodo scietur id quod dicitur eritis enim in aera lo-

14:10 There are, it may be, so many kinds of voices in the quentes 14:10 tam multa ut puta genera linguarum sunt in

world, and none of them is without signification.

14:11 Therefore if I know not the meaning of the voice, mundo et nihil sine uoce est 14:11 si ergo nesciero uirtutem

I shall be unto him that speaketh a barbarian, and he that speaketh shall be a barbarian unto me.

uocis ero ei cui loquor barbarus et qui loquitur mihi barba-

14:12 Even so ye, forasmuch as ye are zealous of spiritual rus 14:12 sic et uos quoniam aemulatores estis spirituum ad

gifts, seek that ye may excel to the edifying of the church.

14:13 Wherefore let him that speaketh in an unknown aedificationem ecclesiae quaerite ut abundetis 14:13 et ideo

tongue pray that he may interpret.

14:14 For if I pray in an unknown tongue, my spirit qui loquitur lingua oret ut interpretetur 14:14 nam si orem

prayeth, but my understanding is unfruitful.

lingua spiritus meus orat mens autem mea sine fructu est

14:15 What is it then? I will pray with the spirit, and I 14:15 quid ergo est orabo spiritu orabo et mente psallam spi-

will pray with the understanding also: I will sing with the spirit, and I will sing with the understanding also.

ritu psallam et mente 14:16 ceterum si benedixeris spiritu

14:16 Else when thou shalt bless with the spirit, how shall he that occupieth the room of the unlearned say Amen at thy giving of thanks, seeing he understandeth not what thou sayest?

qui supplet locum idiotae quomodo dicet amen super tuam

14:17 For thou verily givest thanks well, but the other is benedictionem quoniam quid dicas nescit 14:17 nam tu qui-

not edified.

14:18 I thank my God, I speak with tongues more than ye dem bene gratias agis sed alter non aedificatur 14:18 gratias

all: 14:19 Yet in the church I had rather speak five words with ago Deo quod omnium uestrum lingua loquor 14:19 sed in my understanding, that by my voice I might teach others also, than ten thousand words in an unknown tongue.

ecclesia uolo quinque uerba sensu meo loqui ut et alios in-

14:20 Brethren, be not children in understanding: howbeit struam quam decem milia uerborum in lingua 14:20 fratres in malice be ye children, but in understanding be men.

nolite pueri effici sensibus sed malitia paruuli estote sensibus

14:21 In the law it is written, With men of other tongues autem perfecti estote 14:21 in lege scriptum est quoniam in

and other lips will I speak unto this people; and yet for all that will they not hear me, saith the LORD.

aliis linguis et labiis aliis loquar populo huic et nec sic ex-

14:22 Wherefore tongues are for a sign, not to them that audient me dicit Dominus 14:22 itaque linguae in signum

believe, but to them that believe not: but prophesying serveth not for them that believe not, but for them which

believe. sunt non fidelibus sed infidelibus prophetia autem non infi-14:23 If therefore the whole church be come together into delibus sed fidelibus 14:23 si ergo conueniat uniuersa ecc-

one place, and all speak with tongues, and there come in those that are unlearned, or unbelievers, will they not say that ye are mad?

Epistola Pauli Ad Corinthios I

lesia in unum et omnes linguis loquantur intrent autem idiotae aut infideles nonne dicent quod insanitis 14:24 si au- 14:24 But if all prophesy, and there come in one that be-

lieveth not, or one unlearned, he is convinced of all, he is judged of all:

tem omnes prophetent intret autem quis infidelis uel idiota conuincitur ab omnibus diiudicatur ab omnibus 14:25 oc- 14:25 And thus are the secrets of his heart made manifest;

and so falling down on his face he will worship God, and report that God is in you of a truth.

culta cordis eius manifesta fiunt et ita cadens in faciem adorabit Deum pronuntians quod uere Deus in uobis est 14:26 14:26 How is it then, brethren? when ye come together,

every one of you hath a psalm, hath a doctrine, hath a tongue, hath a revelation, hath an interpretation. Let all

quid ergo est fratres cum conuenitis unusquisque uestrum things be done unto edifying. psalmum habet doctrinam habet apocalypsin habet linguam habet interpretationem habet omnia ad aedificationem fiant 14:27 siue lingua quis loquitur secundum duos aut ut multum 14:27 If any man speak in an unknown tongue, let it be by

two, or at the most by three, and that by course; and let one interpret.

tres et per partes et unus interpretetur 14:28 si autem non

14:28 But if there be no interpreter, let him keep silence in the church; and let him speak to himself, and to God.

fuerit interpres taceat in ecclesia sibi autem loquatur et Deo 14:29 prophetae duo aut tres dicant et ceteri diiudicent 14:30 14:29 Let the prophets speak two or three, and let the other

judge.

14:30 If any thing be revealed to another that sitteth by, let the first hold his peace.

quod si alii reuelatum fuerit sedenti prior taceat 14:31 po-

14:31 For ye may all prophesy one by one, that all may

testis enim omnes per singulos prophetare ut omnes discant _{learn, and all may be comforted.} et omnes exhortentur 14:32 et spiritus prophetarum prophe- 14:32 And the spirits of the prophets are subject to the

prophets.

tis subiecti sunt 14:33 non enim est dissensionis Deus sed 14:33 For God is not the author of confusion, but of peace,

as in all churches of the saints.

pacis sicut in omnibus ecclesiis sanctorum 14:34 mulieres 14:34 Let your women keep silence in the churches: for

it is not permitted unto them to speak; but they are commanded to be under obedience as also saith the law.

in ecclesiis taceant non enim permittitur eis loqui sed subditas esse sicut et lex dicit 14:35 si quid autem uolunt discere 14:35 And if they will learn any thing, let them ask their

husbands at home: for it is a shame for women to speak in the church.

domi uiros suos interrogent turpe est enim mulieri loqui in ecclesia 14:36 an a uobis uerbum Dei processit aut in uos 14:36 What? came the word of God out from you? or

came it unto you only?

solos peruenit 14:37 si quis uidetur propheta esse aut spi- 14:37 If any man think himself to be a prophet, or spiritual,

let him acknowledge that the things that I write unto you are the commandments of the Lord.

ritalis cognoscat quae scribo uobis quia Domini sunt mandata 14:38 si quis autem ignorat ignorabitur 14:39 itaque 14:38 But if any man be ignorant, let him be ignorant.

14:39 Wherefore, brethren, covet to prophesy, and forbid not to speak with tongues.

fratres aemulamini prophetare et loqui linguis nolite prohibere 14:40 omnia autem honeste et secundum ordinem fiant 14:40 Let all things be done decently and in order.

15:1 notum autem uobis facio fratres euangelium quod 15:1 Moreover, brethren, I declare unto you the gospel which I preached unto you, which also ye have received, and wherein ye stand;

praedicaui uobis quod et accepistis in quo et statis 15:2 per

15:2 By which also ye are saved, if ye keep in memory what I preached unto you, unless ye have believed in vain.

quod et saluamini qua ratione praedicauerim uobis si tenetis nisi si frustra credidistis 15:3 tradidi enim uobis in pri- 15:3 For I delivered unto you first of all that which I also

received, how that Christ died for our sins according to the scriptures;

mis quod et accepi quoniam Christus mortuus est pro peccatis nostris secundum scripturas 15:4 et quia sepultus est 15:4 And that he was buried, and that he rose again the

third day according to the scriptures:

Biblia Sacra Vulgata

15:5 And that he was seen of Cephas, then of the twelve: et quia resurrexit tertia die secundum scripturas 15:5 et quia 15:6 After that, he was seen of above five hundred brethren uisus est Cephae et post haec undecim 15:6 deinde uisus est

at once; of whom the greater part remain unto this present, but some are fallen asleep.

plus quam quingentis fratribus simul ex quibus multi manent

15:7 After that, he was seen of James; then of all the apos- usque adhuc quidam autem dormierunt 15:7 deinde uisus tles.

15:8 And last of all he was seen of me also, as of one born est Iacobo deinde apostolis omnibus 15:8 nouissime autem

out of due time.

15:9 For I am the least of the apostles, that am not meet omnium tamquam abortiuo uisus est et mihi 15:9 ego enim

to be called an apostle, because I persecuted the church of God.

sum minimus apostolorum qui non sum dignus uocari apo-

15:10 But by the grace of God I am what I am: and his stolus quoniam persecutus sum ecclesiam Dei 15:10 gra-

grace which was bestowed upon me was not in vain; but I laboured more abundantly than they all: yet not I, but the

grace of God which was with me. tia autem Dei sum id quod sum et gratia eius in me uacua

non fuit sed abundantius illis omnibus laboraui non ego au-

15:11 Therefore whether it were I or they, so we preach, tem sed gratia Dei mecum 15:11 siue enim ego siue illi sic

and so ye believed.

15:12 Now if Christ be preached that he rose from the praedicamus et sic credidistis 15:12 si autem Christus prae-

dead, how say some among you that there is no resurrection of the dead?

dicatur quod resurrexit a mortuis quomodo quidam dicunt

15:13 But if there be no resurrection of the dead, then is in uobis quoniam resurrectio mortuorum non est 15:13 si Christ not risen:

autem resurrectio mortuorum non est neque Christus resur-

15:14 And if Christ be not risen, then is our preaching rexit 15:14 si autem Christus non resurrexit inanis est ergo

vain, and your faith is also vain.

15:15 Yea, and we are found false witnesses of God; be- praedicatio nostra inanis est et fides uestra 15:15 inuenimur

cause we have testified of God that he raised up Christ: whom he raised not up, if so be that the dead rise not.

autem et falsi testes Dei quoniam testimonium diximus aduersus Deum quod suscitauerit Christum quem non suscita-

15:16 For if the dead rise not, then is not Christ raised: uit si mortui non resurgunt 15:16 nam si mortui non resur-

15:17 And if Christ be not raised, your faith is vain; ye are gunt neque Christus resurrexit 15:17 quod si Christus non yet in your sins.

resurrexit uana est fides uestra adhuc enim estis in pecca-

15:18 Then they also which are fallen asleep in Christ are tis uestris 15:18 ergo et qui dormierunt in Christo perierunt

perished.

15:19 If in this life only we have hope in Christ, we are of 15:19 si in hac uita tantum in Christo sperantes sumus mi-

all men most miserable.

15:20 But now is Christ risen from the dead, and become serabiliores sumus omnibus hominibus 15:20 nunc autem

the firstfruits of them that slept.

15:21 For since by man came death, by man came also the Christus resurrexit a mortuis primitiae dormientium 15:21

resurrection of the dead.

quoniam enim per hominem mors et per hominem resurrec-

15:22 For as in Adam all die, even so in Christ shall all be tio mortuorum 15:22 et sicut in Adam omnes moriuntur ita made alive.

15:23 But every man in his own order: Christ the first- et in Christo omnes uiuificabuntur 15:23 unusquisque au-

fruits; afterward they that are Christ’s at his coming.

tem in suo ordine primitiae Christus deinde hii qui sunt Ch-

15:24 Then cometh the end, when he shall have delivered risti in aduentu eius 15:24 deinde finis cum tradiderit reg-

up the kingdom to God, even the Father; when he shall have put down all rule and all authority and power.

num Deo et Patri cum euacuauerit omnem principatum et

15:25 For he must reign, till he hath put all enemies under potestatem et uirtutem 15:25 oportet autem illum regnare

his feet.

Epistola Pauli Ad Corinthios I

donec ponat omnes inimicos sub pedibus eius 15:26 no- 15:26 The last enemy that shall be destroyed is death. uissima autem inimica destruetur mors omnia enim subiecit sub pedibus eius cum autem dicat 15:27 omnia subiecta 15:27 For he hath put all things under his feet. But when

he saith all things are put under him, it is manifest that he is excepted, which did put all things under him.

sunt sine dubio praeter eum qui subiecit ei omnia 15:28

15:28 And when all things shall be subdued unto him, then shall the Son also himself be subject unto him that put all things under him, that God may be all in all.

cum autem subiecta fuerint illi omnia tunc ipse Filius subiectus erit illi qui sibi subiecit omnia ut sit Deus omnia in omnibus 15:29 alioquin quid facient qui baptizantur pro 15:29 Else what shall they do which are baptized for the

dead, if the dead rise not at all? why are they then baptized for the dead?

mortuis si omnino mortui non resurgunt ut quid et baptizantur pro illis 15:30 ut quid et nos periclitamur omni hora 15:30 And why stand we in jeopardy every hour? 15:31 cotidie morior per uestram gloriam fratres quam habeo 15:31 I protest by your rejoicing which I have in Christ

Jesus our LORD, I die daily.

in Christo Iesu Domino nostro 15:32 si secundum homi- 15:32 If after the manner of men I have fought with beasts at Ephesus, what advantageth it me, if the dead rise not? let us eat and drink; for to morrow we die.

nem ad bestias pugnaui Ephesi quid mihi prodeet si mortui non resurgunt manducemus et bibamus cras enim moriemur 15:33 nolite seduci corrumpunt mores bonos conloquia mala 15:33 Be not deceived: evil communications corrupt good

manners.

15:34 euigilate iuste et nolite peccare ignorantiam enim Dei 15:34 Awake to righteousness, and sin not; for some have

not the knowledge of God: I speak this to your shame.

quidam habent ad reuerentiam uobis loquor 15:35 sed dicet 15:35 But some man will say, How are the dead raised up?

and with what body do they come?

aliquis quomodo resurgunt mortui quali autem corpore ueniunt 15:36 insipiens tu quod seminas non uiuificatur nisi 15:36 Thou fool, that which thou sowest is not quickened,

except it die:

prius moriatur 15:37 et quod seminas non corpus quod fu- 15:37 And that which thou sowest, thou sowest not that

body that shall be, but bare grain, it may chance of wheat, or of some other grain:

turum est seminas sed nudum granum ut puta tritici aut alicuius ceterorum 15:38 Deus autem dat illi corpus sicut uo- 15:38 But God giveth it a body as it hath pleased him, and

to every seed his own body.

luit et unicuique seminum proprium corpus 15:39 non om- 15:39 All flesh is not the same flesh: but there is one kind

of flesh of men, another flesh of beasts, another of fishes, and another of birds.

nis caro eadem caro sed alia hominum alia pecorum alia caro uolucrum alia autem piscium 15:40 et corpora caelestia et 15:40 There are also celestial bodies, and bodies terres-

trial: but the glory of the celestial is one, and the glory of the terrestrial is another.

corpora terrestria sed alia quidem caelestium gloria alia autem terrestrium 15:41 alia claritas solis alia claritas lunae et 15:41 There is one glory of the sun, and another glory of

the moon, and another glory of the stars: for one star differeth from another star in glory.

alia claritas stellarum stella enim ab stella differt in claritate 15:42 sic et resurrectio mortuorum seminatur in corruptione 15:42 So also is the resurrection of the dead. It is sown in

corruption; it is raised in incorruption:

surgit in incorruptione 15:43 seminatur in ignobilitate sur- 15:43 It is sown in dishonour; it is raised in glory: it is

sown in weakness; it is raised in power:

git in gloria seminatur in infirmitate surgit in uirtute 15:44 15:44 It is sown a natural body; it is raised a spiritual body.

There is a natural body, and there is a spiritual body.

seminatur corpus animale surgit corpus spiritale si est corpus animale est et spiritale sic et scriptum est 15:45 factus 15:45 And so it is written, The first man Adam was made

a living soul; the last Adam was made a quickening spirit.

est primus homo Adam in animam uiuentem nouissimus Ad-

Biblia Sacra Vulgata

15:46 Howbeit that was not first which is spiritual, but that am in spiritum uiuificantem 15:46 sed non prius quod spiwhich is natural; and afterward that which is spiritual.

15:47 The first man is of the earth, earthy; the second man ritale est sed quod animale est deinde quod spiritale 15:47

is the Lord from heaven.

primus homo de terra terrenus secundus homo de caelo cae-

15:48 As is the earthy, such are they also that are earthy: lestis 15:48 qualis terrenus tales et terreni et qualis caele-

and as is the heavenly, such are they also that are heavenly.

15:49 And as we have borne the image of the earthy, we stis tales et caelestes 15:49 igitur sicut portauimus imagi-

shall also bear the image of the heavenly. 15:50 Now this I say, brethren, that flesh and blood cannot nem terreni portemus et imaginem caelestis 15:50 hoc auinherit the kingdom of God; neither doth corruption inherit incorruption.

tem dico fratres quoniam caro et sanguis regnum Dei possidere non possunt neque corruptio incorruptelam possidebit

15:51 Behold, I shew you a mystery; We shall not all sleep, 15:51 ecce mysterium uobis dico omnes quidem resurgemus but we shall all be changed,

15:52 In a moment, in the twinkling of an eye, at the last sed non omnes inmutabimur 15:52 in momento in ictu oculi

trump: for the trumpet shall sound, and the dead shall be raised incorruptible, and we shall be changed.

in nouissima tuba canet enim et mortui resurgent incorrupti

15:53 For this corruptible must put on incorruption, and et nos inmutabimur 15:53 oportet enim corruptibile hoc in-

this mortal must put on immortality.

duere incorruptelam et mortale hoc induere inmortalitatem

15:54 So when this corruptible shall have put on incorrup- 15:54 cum autem mortale hoc induerit inmortalitatem tunc tion, and this mortal shall have put on immortality, then shall be brought to pass the saying that is written, Death is

swallowed up in victory. fiet sermo qui scriptus est absorta est mors in uictoria 15:55

15:55 O death, where is thy sting? O grave, where is thy victory?

ubi est mors uictoria tua ubi est mors stimulus tuus 15:56

15:56 The sting of death is sin; and the strength of sin is the law.

stimulus autem mortis peccatum est uirtus uero peccati lex

15:57 But thanks be to God, which giveth us the victory 15:57 Deo autem gratias qui dedit nobis uictoriam per Domi-

through our Lord Jesus Christ. 15:58 Therefore, my beloved brethren, be ye stedfast, un- num nostrum Iesum Christum 15:58 itaque fratres mei dimoveable, always abounding in the work of the Lord, forasmuch as ye know that your labour is not in vain in

the Lord. lecti stabiles estote et inmobiles abundantes in opere Domini semper scientes quod labor uester non est inanis in Domino

16:1 Now concerning the collection for the saints, as I 16:1 de collectis autem quae fiunt in sanctos sicut ordinaui

have given order to the churches of Galatia, even so do ye.

ecclesiis Galatiae ita et uos facite 16:2 per unam sabbati

16:2 Upon the first day of the week let every one of you lay by him in store, as God hath prospered him, that there be no gatherings when I come.

unusquisque uestrum apud se ponat recondens quod ei bene-

16:3 And when I come, whomsoever ye shall approve by placuerit ut non cum uenero tunc collectae fiant 16:3 cum

your letters, them will I send to bring your liberality unto Jerusalem.

autem praesens fuero quos probaueritis per epistulas hos

16:4 And if it be meet that I go also, they shall go with me. mittam perferre gratiam uestram in Hierusalem 16:4 quod

16:5 Now I will come unto you, when I shall pass through si dignum fuerit ut et ego eam mecum ibunt 16:5 ueniam Macedonia: for I do pass through Macedonia.

autem ad uos cum Macedoniam pertransiero nam Macedo-

16:6 And it may be that I will abide, yea, and winter with niam pertransibo 16:6 apud uos autem forsitan manebo uel

you, that ye may bring me on my journey whithersoever I go.

etiam hiemabo ut uos me deducatis quocumque iero 16:7

16:7 For I will not see you now by the way; but I trust to tarry a while with you, if the Lord permit.

nolo enim uos modo in transitu uidere spero enim me aliquantum temporis manere apud uos si Dominus permiserit

Epistola Pauli Ad Corinthios I

16:8 permanebo autem Ephesi usque ad pentecosten 16:9 16:8 But I will tarry at Ephesus until Pentecost.

16:9 For a great door and effectual is opened unto me, and there are many adversaries.

ostium enim mihi apertum est magnum et euidens et aduersarii multi 16:10 si autem uenerit Timotheus uidete ut 16:10 Now if Timotheus come, see that he may be with

you without fear: for he worketh the work of the Lord, as I also do.

sine timore sit apud uos opus enim Domini operatur sicut et ego 16:11 ne quis ergo illum spernat deducite autem il- 16:11 Let no man therefore despise him: but conduct him

forth in peace, that he may come unto me: for I look for him with the brethren.

lum in pace ut ueniat ad me expecto enim illum cum frat-

ribus 16:12 de Apollo autem fratre multum rogaui eum ut 16:12 As touching our brother Apollos, I greatly desired him to come unto you with the brethren: but his will was not at all to come at this time; but he will come when he

ueniret ad uos cum fratribus et utique non fuit uoluntas ut shall have convenient time. nunc ueniret ueniet autem cum ei uacuum fuerit 16:13 uigi- 16:13 Watch ye, stand fast in the faith, quit you like men,

be strong.

late state in fide uiriliter agite et confortamini 16:14 omnia 16:14 Let all your things be done with charity. uestra in caritate fiant 16:15 obsecro autem uos fratres no- 16:15 I beseech you, brethren, (ye know the house of

Stephanas, that it is the firstfruits of Achaia, and that they have addicted themselves to the ministry of the saints,)

stis domum Stephanae et Fortunati quoniam sunt primitiae Achaiae et in ministerium sanctorum ordinauerunt se ipsos 16:16 ut et uos subditi sitis eiusmodi et omni cooperanti et la- 16:16 That ye submit yourselves unto such, and to every

one that helpeth with us, and laboureth.

boranti 16:17 gaudeo autem in praesentia Stephanae et For- 16:17 I am glad of the coming of Stephanas and Fortunatus

and Achaicus: for that which was lacking on your part they have supplied.

tunati et Achaici quoniam id quod uobis deerat ipsi supple-

uerunt 16:18 refecerunt enim et meum spiritum et uestrum 16:18 For they have refreshed my spirit and your’s: there-

fore acknowledge ye them that are such.

cognoscite ergo qui eiusmodi sunt 16:19 salutant uos eccle- 16:19 The churches of Asia salute you. Aquila and

Priscilla salute you much in the Lord, with the church that is in their house.

siae Asiae salutant uos in Domino multum Aquila et Prisca cum domestica sua ecclesia 16:20 salutant uos fratres om- 16:20 All the brethren greet you. Greet ye one another

with an holy kiss.

nes salutate inuicem in osculo sancto 16:21 salutatio mea 16:21 The salutation of me Paul with mine own hand. manu Pauli 16:22 si quis non amat Dominum Iesum Chris- 16:22 If any man love not the Lord Jesus Christ, let him

be Anathema Maranatha.

tum sit anathema maranatha 16:23 gratia Domini Iesu uobi- 16:23 The grace of our Lord Jesus Christ be with you. scum 16:24 caritas mea cum omnibus uobis in Christo Iesu 16:24 My love be with you all in Christ Jesus. Amen. amen

Biblia Sacra Vulgata

MDCLIV

Epistola Pauli Ad Corinthios II

1:1 Paulus apostolus Iesu Christi per uoluntatem Dei et Ti- 1:1 Paul, an apostle of Jesus Christ by the will of God, and

Timothy our brother, unto the church of God which is at Corinth, with all the saints which are in all Achaia:

motheus frater ecclesiae Dei quae est Corinthi cum sanctis omnibus qui sunt in uniuersa Achaia 1:2 gratia uobis et pax 1:2 Grace be to you and peace from God our Father, and

from the Lord Jesus Christ.

a Deo Patre nostro et Domino Iesu Christo 1:3 benedictus 1:3 Blessed be God, even the Father of our Lord Jesus

Christ, the Father of mercies, and the God of all comfort;

Deus et Pater Domini nostri Iesu Christi Pater misericordiarum et Deus totius consolationis 1:4 qui consolatur nos in 1:4 Who comforteth us in all our tribulation, that we may

be able to comfort them which are in any trouble, by the comfort wherewith we ourselves are comforted of God.

omni tribulatione nostra ut possimus et ipsi consolari eos qui in omni pressura sunt per exhortationem qua exhortamur et ipsi a Deo 1:5 quoniam sicut abundant passiones Christi 1:5 For as the sufferings of Christ abound in us, so our

consolation also aboundeth by Christ.

in nobis ita et per Christum abundat consolatio nostra 1:6 1:6 And whether we be afflicted, it is for your consola-

tion and salvation, which is effectual in the enduring of the same sufferings which we also suffer: or whether we

siue autem tribulamur pro uestra exhortatione et salute siue be comforted, it is for your consolation and salvation. exhortamur pro uestra exhortatione quae operatur in tolerantia earundem passionum quas et nos patimur 1:7 et spes 1:7 And our hope of you is stedfast, knowing, that as ye

are partakers of the sufferings, so shall ye be also of the consolation.

nostra firma pro uobis scientes quoniam sicut socii passionum estis sic eritis et consolationis 1:8 non enim uolumus 1:8 For we would not, brethren, have you ignorant of our

trouble which came to us in Asia, that we were pressed out of measure, above strength, insomuch that we despaired

ignorare uos fratres de tribulatione nostra quae facta est in even of life: Asia quoniam supra modum grauati sumus supra uirtutem ita ut taederet nos etiam uiuere 1:9 sed ipsi in nobis ipsis 1:9 But we had the sentence of death in ourselves, that we

should not trust in ourselves, but in God which raiseth the dead:

responsum mortis habuimus ut non simus fidentes in nobis sed in Deo qui suscitat mortuos 1:10 qui de tantis pericu- 1:10 Who delivered us from so great a death, and doth

deliver: in whom we trust that he will yet deliver us;

lis eripuit nos et eruet in quem speramus quoniam et adhuc eripiet 1:11 adiuuantibus et uobis in oratione pro nobis ut 1:11 Ye also helping together by prayer for us, that for

the gift bestowed upon us by the means of many persons thanks may be given by many on our behalf.

ex multis personis eius quae in nobis est donationis per multos gratiae agantur pro nobis 1:12 nam gloria nostra haec 1:12 For our rejoicing is this, the testimony of our con-

science, that in simplicity and godly sincerity, not with fleshly wisdom, but by the grace of God, we have had our conversation in the world, and more abundantly to youward.

est testimonium conscientiae nostrae quod in simplicitate et sinceritate Dei et non in sapientia carnali sed in gratia Dei conuersati sumus in mundo abundantius autem ad uos 1:13 1:13 For we write none other things unto you, than what

ye read or acknowledge; and I trust ye shall acknowledge even to the end;

non enim alia scribimus uobis quam quae legistis et cognoscitis spero autem quod usque in finem cognoscetis 1:14 1:14 As also ye have acknowledged us in part, that we are

your rejoicing, even as ye also are our’s in the day of the Lord Jesus.

sicut et cognouistis nos ex parte quia gloria uestra sumus sicut et uos nostra in die Domini nostri Iesu Christi 1:15 et 1:15 And in this confidence I was minded to come unto

you before, that ye might have a second benefit;

hac confidentia uolui prius uenire ad uos ut secundam gratiam haberetis 1:16 et per uos transire in Macedoniam et 1:16 And to pass by you into Macedonia, and to come

again out of Macedonia unto you, and of you to be brought on my way toward Judaea.

Biblia Sacra Vulgata

iterum a Macedonia uenire ad uos et a uobis deduci in Iu-

1:17 When I therefore was thus minded, did I use light- daeam 1:17 cum hoc ergo uoluissem numquid leuitate usus

ness? or the things that I purpose, do I purpose according to the flesh, that with me there should be yea yea, and nay

nay? sum aut quae cogito secundum carnem cogito ut sit apud me 1:18 But as God is true, our word toward you was not yea est et non 1:18 fidelis autem Deus quia sermo noster qui

and nay. 1:19 For the Son of God, Jesus Christ, who was preached fit apud uos non est in illo est et non 1:19 Dei enim Fiamong you by us, even by me and Silvanus and Timotheus, was not yea and nay, but in him was yea.

lius Iesus Christus qui in uobis per nos praedicatus est per me et Siluanum et Timotheum non fuit est et non sed est in

1:20 For all the promises of God in him are yea, and in illo fuit 1:20 quotquot enim promissiones Dei sunt in illo

him Amen, unto the glory of God by us. 1:21 Now he which stablisheth us with you in Christ, and est ideo et per ipsum amen Deo ad gloriam nostram 1:21 hath anointed us, is God;

qui autem confirmat nos uobiscum in Christum et qui unxit

1:22 Who hath also sealed us, and given the earnest of the nos Deus 1:22 et qui signauit nos et dedit pignus Spiritus

Spirit in our hearts.

1:23 Moreover I call God for a record upon my soul, that in cordibus nostris 1:23 ego autem testem Deum inuoco in

to spare you I came not as yet unto Corinth.

animam meam quod parcens uobis non ueni ultra Corinthum

1:24 Not for that we have dominion over your faith, but 1:24 non quia dominamur fidei uestrae sed adiutores sumus

are helpers of your joy: for by faith ye stand.

gaudii uestri nam fide stetistis

2:1 But I determined this with myself, that I would not 2:1 statui autem hoc ipse apud me ne iterum in tristitia ue-

come again to you in heaviness.

2:2 For if I make you sorry, who is he then that maketh me nirem ad uos 2:2 si enim ego contristo uos et quis est qui

glad, but the same which is made sorry by me?

2:3 And I wrote this same unto you, lest, when I came, I me laetificet nisi qui contristatur ex me 2:3 et hoc ipsum

should have sorrow from them of whom I ought to rejoice; having confidence in you all, that my joy is the joy of you

all. scripsi ut non cum uenero tristitiam super tristitiam habeam

de quibus oportuerat me gaudere confidens in omnibus uo-

2:4 For out of much affliction and anguish of heart I wrote bis quia meum gaudium omnium uestrum est 2:4 nam ex unto you with many tears; not that ye should be grieved, but that ye might know the love which I have more abun-

dantly unto you. multa tribulatione et angustia cordis scripsi uobis per multas

lacrimas non ut contristemini sed ut sciatis quam caritatem

2:5 But if any have caused grief, he hath not grieved me, habeo abundantius in uobis 2:5 si quis autem contristauit

but in part: that I may not overcharge you all.

non me contristauit sed ex parte ut non onerem omnes uos

2:6 Sufficient to such a man is this punishment, which was 2:6 sufficit illi qui eiusmodi est obiurgatio haec quae fit a

inflicted of many. 2:7 So that contrariwise ye ought rather to forgive him, and pluribus 2:7 ita ut e contra magis donetis et consolemini ne comfort him, lest perhaps such a one should be swallowed up with overmuch sorrow.

forte abundantiori tristitia absorbeatur qui eiusmodi est 2:8

2:8 Wherefore I beseech you that ye would confirm your love toward him.

propter quod obsecro uos ut confirmetis in illum caritatem

2:9 For to this end also did I write, that I might know the 2:9 ideo enim et scripsi ut cognoscam experimentum uestrum

proof of you, whether ye be obedient in all things.

2:10 To whom ye forgive any thing, I forgive also: for if an in omnibus oboedientes sitis 2:10 cui autem aliquid do-

I forgave any thing, to whom I forgave it, for your sakes forgave I it in the person of Christ;

natis et ego nam et ego quod donaui si quid donaui propter

2:11 Lest Satan should get an advantage of us: for we are uos in persona Christi 2:11 ut non circumueniamur a Satana

not ignorant of his devices.

Epistola Pauli Ad Corinthios II

non enim ignoramus cogitationes eius 2:12 cum uenissem 2:12 Furthermore, when I came to Troas to preach Christ’s

gospel, and a door was opened unto me of the Lord,

autem Troadem propter euangelium Christi et ostium mihi apertum esset in Domino 2:13 non habui requiem spiritui 2:13 I had no rest in my spirit, because I found not Ti-

tus my brother: but taking my leave of them, I went from thence into Macedonia.

meo eo quod non inuenerim Titum fratrem meum sed ualefaciens eis profectus sum in Macedoniam 2:14 Deo autem 2:14 Now thanks be unto God, which always causeth us to

triumph in Christ, and maketh manifest the savour of his knowledge by us in every place.

gratias qui semper triumphat nos in Christo Iesu et odorem notitiae suae manifestat per nos in omni loco 2:15 quia Ch- 2:15 For we are unto God a sweet savour of Christ, in them

that are saved, and in them that perish:

risti bonus odor sumus Deo in his qui salui fiunt et in his qui pereunt 2:16 aliis quidem odor mortis in mortem aliis 2:16 To the one we are the savour of death unto death;

and to the other the savour of life unto life. And who is sufficient for these things?

autem odor uitae in uitam et ad haec quis tam idoneus 2:17

2:17 For we are not as many, which corrupt the word of God: but as of sincerity, but as of God, in the sight of God speak we in Christ.

non enim sumus sicut plurimi adulterantes uerbum Dei sed ex sinceritate sed sicut ex Deo coram Deo in Christo loquimur

3:1 incipimus iterum nosmet ipsos commendare aut num- 3:1 Do we begin again to commend ourselves? or need

we, as some others, epistles of commendation to you, or letters of commendation from you?

quid egemus sicut quidam commendaticiis epistulis ad uos aut ex uobis 3:2 epistula nostra uos estis scripta in cordi- 3:2 Ye are our epistle written in our hearts, known and read

of all men:

bus nostris quae scitur et legitur ab omnibus hominibus 3:3 3:3 Forasmuch as ye are manifestly declared to be the epis-

tle of Christ ministered by us, written not with ink, but with the Spirit of the living God; not in tables of stone, but

manifestati quoniam epistula estis Christi ministrata a nobis in fleshy tables of the heart. et scripta non atramento sed Spiritu Dei uiui non in tabulis lapideis sed in tabulis cordis carnalibus 3:4 fiduciam autem 3:4 And such trust have we through Christ to God-ward: talem habemus per Christum ad Deum 3:5 non quod suffi- 3:5 Not that we are sufficient of ourselves to think any

thing as of ourselves; but our sufficiency is of God;

cientes simus cogitare aliquid a nobis quasi ex nobis sed sufficientia nostra ex Deo est 3:6 qui et idoneos nos fecit mi- 3:6 Who also hath made us able ministers of the new tes-

tament; not of the letter, but of the spirit: for the letter killeth, but the spirit giveth life.

nistros noui testamenti non litterae sed Spiritus littera enim occidit Spiritus autem uiuificat 3:7 quod si ministratio mor- 3:7 But if the ministration of death, written and engraven

in stones, was glorious, so that the children of Israel could not stedfastly behold the face of Moses for the glory of his

tis litteris deformata in lapidibus fuit in gloria ita ut non pos- countenance; which glory was to be done away: sent intendere filii Israhel in faciem Mosi propter gloriam

uultus eius quae euacuatur 3:8 quomodo non magis mini- 3:8 How shall not the ministration of the spirit be rather

glorious?

stratio Spiritus erit in gloria 3:9 nam si ministratio damna- 3:9 For if the ministration of condemnation be glory, much

more doth the ministration of righteousness exceed in glory.

tionis gloria est multo magis abundat ministerium iustitiae in gloria 3:10 nam nec glorificatum est quod claruit in hac 3:10 For even that which was made glorious had no glory

in this respect, by reason of the glory that excelleth.

parte propter excellentem gloriam 3:11 si enim quod eua- 3:11 For if that which is done away was glorious, much

more that which remaineth is glorious.

cuatur per gloriam est multo magis quod manet in gloria est

Biblia Sacra Vulgata

3:12 Seeing then that we have such hope, we use great 3:12 habentes igitur talem spem multa fiducia utimur 3:13

plainness of speech: 3:13 And not as Moses, which put a vail over his face, that the children of Israel could not stedfastly look to the end of that which is abolished:

et non sicut Moses ponebat uelamen super faciem suam ut non intenderent filii Israhel in faciem eius quod euacuatur

3:14 But their minds were blinded: for until this day re- 3:14 sed obtusi sunt sensus eorum usque in hodiernum enim

maineth the same vail untaken away in the reading of the old testament; which vail is done away in Christ.

diem id ipsum uelamen in lectione ueteris testamenti ma-

3:15 But even unto this day, when Moses is read, the vail net non reuelatum quoniam in Christo euacuatur 3:15 sed

is upon their heart.

usque in hodiernum diem cum legitur Moses uelamen est

3:16 Nevertheless when it shall turn to the Lord, the vail positum super cor eorum 3:16 cum autem conuersus fueshall be taken away. 3:17 Now the Lord is that Spirit: and where the Spirit of rit ad Deum aufertur uelamen 3:17 Dominus autem Spirithe Lord is, there is liberty.

3:18 But we all, with open face beholding as in a glass the tus est ubi autem Spiritus Domini ibi libertas 3:18 nos uero

glory of the Lord, are changed into the same image from glory to glory, even as by the Spirit of the LORD.

omnes reuelata facie gloriam Domini speculantes in eandem imaginem transformamur a claritate in claritatem tamquam a Domini Spiritu

4:1 Therefore seeing we have this ministry, as we have 4:1 ideo habentes hanc ministrationem iuxta quod miseri-

received mercy, we faint not; 4:2 But have renounced the hidden things of dishonesty, cordiam consecuti sumus non deficimus 4:2 sed abdicamus not walking in craftiness, nor handling the word of God deceitfully; but by manifestation of the truth commending

ourselves to every man’s conscience in the sight of God. occulta dedecoris non ambulantes in astutia neque adulterantes uerbum Dei sed in manifestatione ueritatis commendantes nosmet ipsos ad omnem conscientiam hominum coram

4:3 But if our gospel be hid, it is hid to them that are lost: Deo 4:3 quod si etiam opertum est euangelium nostrum 4:4 In whom the god of this world hath blinded the minds in his qui pereunt est opertum 4:4 in quibus deus huius of them which believe not, lest the light of the glorious gospel of Christ, who is the image of God, should shine

unto them. saeculi excaecauit mentes infidelium ut non fulgeat inlumi-

4:5 For we preach not ourselves, but Christ Jesus the Lord; natio euangelii gloriae Christi qui est imago Dei 4:5 non and ourselves your servants for Jesus’ sake.

enim nosmet ipsos praedicamus sed Iesum Christum Domi-

4:6 For God, who commanded the light to shine out of num nos autem seruos uestros per Iesum 4:6 quoniam Deus

darkness, hath shined in our hearts, to give the light of the knowledge of the glory of God in the face of Jesus Christ.

qui dixit de tenebris lucem splendescere qui inluxit in cordibus nostris ad inluminationem scientiae claritatis Dei in

4:7 But we have this treasure in earthen vessels, that the facie Christi Iesu 4:7 habemus autem thesaurum istum in excellency of the power may be of God, and not of us.

uasis fictilibus ut sublimitas sit uirtutis Dei et non ex nobis

4:8 We are troubled on every side, yet not distressed; we 4:8 in omnibus tribulationem patimur sed non angustiamur

are perplexed, but not in despair;

4:9 Persecuted, but not forsaken; cast down, but not de- aporiamur sed non destituimur 4:9 persecutionem patimur

stroyed;

4:10 Always bearing about in the body the dying of the sed non derelinquimur deicimur sed non perimus 4:10 sem-

Lord Jesus, that the life also of Jesus might be made manifest in our body.

per mortificationem Iesu in corpore nostro circumferentes ut

4:11 For we which live are alway delivered unto death for et uita Iesu in corporibus nostris manifestetur 4:11 semper

Jesus’ sake, that the life also of Jesus might be made manifest in our mortal flesh.

Epistola Pauli Ad Corinthios II

enim nos qui uiuimus in mortem tradimur propter Iesum ut et uita Iesu manifestetur in carne nostra mortali 4:12 ergo 4:12 So then death worketh in us, but life in you.

mors in nobis operatur uita autem in uobis 4:13 habentes 4:13 We having the same spirit of faith, according as it is written, I believed, and therefore have I spoken; we also believe, and therefore speak;

autem eundem spiritum fidei sicut scriptum est credidi propter quod locutus sum et nos credimus propter quod et loquimur 4:14 scientes quoniam qui suscitauit Iesum et nos cum 4:14 Knowing that he which raised up the Lord Jesus shall

raise up us also by Jesus, and shall present us with you.

Iesu suscitabit et constituet uobiscum 4:15 omnia enim pro- 4:15 For all things are for your sakes, that the abundant

grace might through the thanksgiving of many redound to the glory of God.

pter uos ut gratia abundans per multos gratiarum actione

abundet in gloriam Dei 4:16 propter quod non deficimus 4:16 For which cause we faint not; but though our outward

man perish, yet the inward man is renewed day by day.

sed licet is qui foris est noster homo corrumpitur tamen is qui intus est renouatur de die in diem 4:17 id enim quod 4:17 For our light affliction, which is but for a moment,

worketh for us a far more exceeding and eternal weight of glory;

in praesenti est momentaneum et leue tribulationis nostrae supra modum in sublimitatem aeternum gloriae pondus operatur nobis 4:18 non contemplantibus nobis quae uidentur 4:18 While we look not at the things which are seen, but

at the things which are not seen: for the things which are seen are temporal; but the things which are not seen are

sed quae non uidentur quae enim uidentur temporalia sunt eternal. quae autem non uidentur aeterna sunt

5:1 scimus enim quoniam si terrestris domus nostra huius 5:1 For we know that if our earthly house of this tabernacle were dissolved, we have a building of God, an house not made with hands, eternal in the heavens.

habitationis dissoluatur quod aedificationem ex Deo habea-

mus domum non manufactam aeternam in caelis 5:2 nam 5:2 For in this we groan, earnestly desiring to be clothed

upon with our house which is from heaven:

et in hoc ingemescimus habitationem nostram quae de caelo est superindui cupientes 5:3 si tamen uestiti non nudi in- 5:3 If so be that being clothed we shall not be found naked.

ueniamur 5:4 nam et qui sumus in tabernaculo ingemesci- 5:4 For we that are in this tabernacle do groan, being burdened: not for that we would be unclothed, but clothed upon, that mortality might be swallowed up of life.

mus grauati eo quod nolumus expoliari sed superuestiri ut absorbeatur quod mortale est a uita 5:5 qui autem efficit 5:5 Now he that hath wrought us for the selfsame thing is

God, who also hath given unto us the earnest of the Spirit.

nos in hoc ipsum Deus qui dedit nobis pignus Spiritus 5:6 5:6 Therefore we are always confident, knowing that,

whilst we are at home in the body, we are absent from the Lord:

audentes igitur semper et scientes quoniam dum sumus in corpore peregrinamur a Domino 5:7 per fidem enim am- 5:7 (For we walk by faith, not by sight:)

bulamus et non per speciem 5:8 audemus autem et bonam 5:8 We are confident, I say, and willing rather to be absent

from the body, and to be present with the Lord.

uoluntatem habemus magis peregrinari a corpore et praesen-

tes esse ad Deum 5:9 et ideo contendimus siue absentes 5:9 Wherefore we labour, that, whether present or absent,

we may be accepted of him.

siue praesentes placere illi 5:10 omnes enim nos manifes- 5:10 For we must all appear before the judgment seat of

Christ; that every one may receive the things done in his body, according to that he hath done, whether it be good

tari oportet ante tribunal Christi ut referat unusquisque prop- or bad. ria corporis prout gessit siue bonum siue malum 5:11 scien- 5:11 Knowing therefore the terror of the Lord, we per-

suade men; but we are made manifest unto God; and I trust also are made manifest in your consciences.

Biblia Sacra Vulgata

tes ergo timorem Domini hominibus suademus Deo autem manifesti sumus spero autem et in conscientiis uestris ma-

5:12 For we commend not ourselves again unto you, but nifestos nos esse 5:12 non iterum nos commendamus uobis

give you occasion to glory on our behalf, that ye may have somewhat to answer them which glory in appearance, and

not in heart. sed occasionem damus uobis gloriandi pro nobis ut habeatis 5:13 For whether we be beside ourselves, it is to God: or ad eos qui in facie gloriantur et non in corde 5:13 siue enim

whether we be sober, it is for your cause.

5:14 For the love of Christ constraineth us; because we mente excedimus Deo siue sobrii sumus uobis 5:14 caritas

thus judge, that if one died for all, then were all dead:

enim Christi urget nos aestimantes hoc quoniam si unus pro

5:15 And that he died for all, that they which live should omnibus mortuus est ergo omnes mortui sunt 5:15 et pro not henceforth live unto themselves, but unto him which died for them, and rose again.

omnibus mortuus est ut et qui uiuunt iam non sibi uiuant

5:16 Wherefore henceforth know we no man after the sed ei qui pro ipsis mortuus est et resurrexit 5:16 itaque nos

flesh: yea, though we have known Christ after the flesh, yet now henceforth know we him no more.

ex hoc neminem nouimus secundum carnem et si cognouimus secundum carnem Christum sed nunc iam non nouimus

5:17 Therefore if any man be in Christ, he is a new crea- 5:17 si qua ergo in Christo noua creatura uetera transierunt

ture: old things are passed away; behold, all things are become new.

ecce facta sunt noua 5:18 omnia autem ex Deo qui recon-

5:18 And all things are of God, who hath reconciled us to himself by Jesus Christ, and hath given to us the ministry of reconciliation;

ciliauit nos sibi per Christum et dedit nobis ministerium re-

5:19 To wit, that God was in Christ, reconciling the world conciliationis 5:19 quoniam quidem Deus erat in Christo unto himself, not imputing their trespasses unto them; and hath committed unto us the word of reconciliation.

mundum reconcilians sibi non reputans illis delicta ipsorum

5:20 Now then we are ambassadors for Christ, as though et posuit in nobis uerbum reconciliationis 5:20 pro Christo

God did beseech you by us: we pray you in Christ’s stead, be ye reconciled to God.

ergo legationem fungimur tamquam Deo exhortante per nos

5:21 For he hath made him to be sin for us, who knew no obsecramus pro Christo reconciliamini Deo 5:21 eum qui

sin; that we might be made the righteousness of God in him.

non nouerat peccatum pro nobis peccatum fecit ut nos efficeremur iustitia Dei in ipso

6:1 We then, as workers together with him, beseech you 6:1 adiuuantes autem et exhortamur ne in uacuum gratiam

also that ye receive not the grace of God in vain.

6:2 (For he saith, I have heard thee in a time accepted, and Dei recipiatis 6:2 ait enim tempore accepto exaudiui te et

in the day of salvation have I succoured thee: behold, now is the accepted time; behold, now is the day of salvation.)

in die salutis adiuuaui te ecce nunc tempus acceptabile ecce

6:3 Giving no offence in any thing, that the ministry be not nunc dies salutis 6:3 nemini dantes ullam offensionem ut blamed: 6:4 But in all things approving ourselves as the ministers non uituperetur ministerium 6:4 sed in omnibus exhibeaof God, in much patience, in afflictions, in necessities, in distresses,

mus nosmet ipsos sicut Dei ministros in multa patientia in

6:5 In stripes, in imprisonments, in tumults, in labours, in tribulationibus in necessitatibus in angustiis 6:5 in plagis

watchings, in fastings;

in carceribus in seditionibus in laboribus in uigiliis in ieiu-

6:6 By pureness, by knowledge, by longsuffering, by kind- niis 6:6 in castitate in scientia in longanimitate in suaui-

ness, by the Holy Ghost, by love unfeigned,

6:7 By the word of truth, by the power of God, by the tate in Spiritu Sancto in caritate non ficta 6:7 in uerbo ue-

armour of righteousness on the right hand and on the left,

ritatis in uirtute Dei per arma iustitiae a dextris et sinistris

Epistola Pauli Ad Corinthios II

6:8 per gloriam et ignobilitatem per infamiam et bonam fa- 6:8 By honour and dishonour, by evil report and good re-

port: as deceivers, and yet true;

mam ut seductores et ueraces sicut qui ignoti et cogniti 6:9 6:9 As unknown, and yet well known; as dying, and, be-

hold, we live; as chastened, and not killed;

quasi morientes et ecce uiuimus ut castigati et non mortificati 6:10 quasi tristes semper autem gaudentes sicut egentes 6:10 As sorrowful, yet alway rejoicing; as poor, yet mak-

ing many rich; as having nothing, and yet possessing all things.

multos autem locupletantes tamquam nihil habentes et omnia possidentes 6:11 os nostrum patet ad uos o Corinthii cor 6:11 O ye Corinthians, our mouth is open unto you, our

heart is enlarged.

nostrum dilatatum est 6:12 non angustiamini in nobis angu- 6:12 Ye are not straitened in us, but ye are straitened in

your own bowels.

stiamini autem in uisceribus uestris 6:13 eandem autem ha- 6:13 Now for a recompence in the same, (I speak as unto

my children,) be ye also enlarged.

bentes remunerationem tamquam filiis dico dilatamini et uos 6:14 nolite iugum ducere cum infidelibus quae enim partici- 6:14 Be ye not unequally yoked together with unbelievers:

for what fellowship hath righteousness with unrighteousness? and what communion hath light with darkness?

patio iustitiae cum iniquitate aut quae societas luci ad tenebras 6:15 quae autem conuentio Christi ad Belial aut quae 6:15 And what concord hath Christ with Belial? or what

part hath he that believeth with an infidel?

pars fideli cum infidele 6:16 qui autem consensus templo 6:16 And what agreement hath the temple of God with

idols? for ye are the temple of the living God; as God hath said, I will dwell in them, and walk in them; and I will be

Dei cum idolis uos enim estis templum Dei uiui sicut dicit their God, and they shall be my people. Deus quoniam inhabitabo in illis et inambulabo et ero illorum Deus et ipsi erunt mihi populus 6:17 propter quod exite 6:17 Wherefore come out from among them, and be ye

separate, saith the Lord, and touch not the unclean thing; and I will receive you.

de medio eorum et separamini dicit Dominus et inmundum ne tetigeritis 6:18 et ego recipiam uos et ero uobis in patrem 6:18 And will be a Father unto you, and ye shall be my

sons and daughters, saith the Lord Almighty.

et uos eritis mihi in filios et filias dicit Dominus omnipotens

7:1 has igitur habentes promissiones carissimi munde- 7:1 Having therefore these promises, dearly beloved, let us cleanse ourselves from all filthiness of the flesh and spirit, perfecting holiness in the fear of God.

mus nos ab omni inquinamento carnis et spiritus perficientes sanctificationem in timore Dei 7:2 capite nos neminem lae- 7:2 Receive us; we have wronged no man, we have cor-

rupted no man, we have defrauded no man.

simus neminem corrupimus neminem circumuenimus 7:3 7:3 I speak not this to condemn you: for I have said before,

that ye are in our hearts to die and live with you.

non ad condemnationem dico praedixi enim quod in cordibus nostris estis ad conmoriendum et ad conuiuendum 7:4 7:4 Great is my boldness of speech toward you, great is my

glorying of you: I am filled with comfort, I am exceeding joyful in all our tribulation.

multa mihi fiducia est apud uos multa mihi gloriatio pro uobis repletus sum consolatione superabundo gaudio in omni tribulatione nostra 7:5 nam et cum uenissemus Macedo- 7:5 For, when we were come into Macedonia, our flesh

had no rest, but we were troubled on every side; without were fightings, within were fears.

niam nullam requiem habuit caro nostra sed omnem tribulationem passi foris pugnae intus timores 7:6 sed qui con- 7:6 Nevertheless God, that comforteth those that are cast

down, comforted us by the coming of Titus;

solatur humiles consolatus est nos Deus in aduentu Titi 7:7 7:7 And not by his coming only, but by the consolation

wherewith he was comforted in you, when he told us your earnest desire, your mourning, your fervent mind toward

non solum autem in aduentu eius sed etiam in solacio quo me; so that I rejoiced the more. consolatus est in uobis referens nobis uestrum desiderium

Biblia Sacra Vulgata

uestrum fletum uestram aemulationem pro me ita ut magis

7:8 For though I made you sorry with a letter, I do not gauderem 7:8 quoniam et si contristaui uos in epistula non

repent, though I did repent: for I perceive that the same epistle hath made you sorry, though it were but for a sea-

son. me paenitet et si paeniteret uidens quod epistula illa et si ad 7:9 Now I rejoice, not that ye were made sorry, but that horam uos contristauit 7:9 nunc gaudeo non quia contris-

ye sorrowed to repentance: for ye were made sorry after a godly manner, that ye might receive damage by us in

nothing. tati estis sed quia contristati estis ad paenitentiam contristati

enim estis secundum Deum ut in nullo detrimentum patia-

7:10 For godly sorrow worketh repentance to salvation not mini ex nobis 7:10 quae enim secundum Deum tristitia est to be repented of: but the sorrow of the world worketh death.

paenitentiam in salutem stabilem operatur saeculi autem tri-

7:11 For behold this selfsame thing, that ye sorrowed after stitia mortem operatur 7:11 ecce enim hoc ipsum secundum a godly sort, what carefulness it wrought in you, yea, what clearing of yourselves, yea, what indignation, yea, what fear, yea, what vehement desire, yea, what zeal, yea, what revenge! In all things ye have approved yourselves to be clear in this matter.

Deum contristari uos quantam in uobis operatur sollicitudinem sed defensionem sed indignationem sed timorem sed desiderium sed aemulationem sed uindictam in omnibus ex-

7:12 Wherefore, though I wrote unto you, I did it not for hibuistis uos incontaminatos esse negotio 7:12 igitur et si

his cause that had done the wrong, nor for his cause that suffered wrong, but that our care for you in the sight of

God might appear unto you. scripsi uobis non propter eum qui fecit iniuriam nec prop-

ter eum qui passus est sed ad manifestandam sollicitudinem

7:13 Therefore we were comforted in your comfort: yea, nostram quam pro uobis habemus ad uos coram Deo 7:13 and exceedingly the more joyed we for the joy of Titus, because his spirit was refreshed by you all.

ideo consolati sumus in consolatione autem nostra abundantius magis gauisi sumus super gaudium Titi quia refectus est

7:14 For if I have boasted any thing to him of you, I am spiritus eius ab omnibus uobis 7:14 et si quid apud illum de

not ashamed; but as we spake all things to you in truth, even so our boasting, which I made before Titus, is found

a truth. uobis gloriatus sum non sum confusus sed sicut omnia uobis

in ueritate locuti sumus ita et gloriatio nostra quae fuit ad Ti-

7:15 And his inward affection is more abundant toward tum ueritas facta est 7:15 et uiscera eius abundantius in uos

you, whilst he remembereth the obedience of you all, how with fear and trembling ye received him.

sunt reminiscentis omnium uestrum oboedientiam quomodo

7:16 I rejoice therefore that I have confidence in you in all cum timore et tremore excepistis eum 7:16 gaudeo quod in things.

omnibus confido in uobis

8:1 Moreover, brethren, we do you to wit of the grace of 8:1 notam autem facimus uobis fratres gratiam Dei quae

God bestowed on the churches of Macedonia;

8:2 How that in a great trial of affliction the abundance of data est in ecclesiis Macedoniae 8:2 quod in multo experi-

their joy and their deep poverty abounded unto the riches of their liberality.

mento tribulationis abundantia gaudii ipsorum et altissima paupertas eorum abundauit in diuitias simplicitatis eorum

8:3 For to their power, I bear record, yea, and beyond their 8:3 quia secundum uirtutem testimonium illis reddo et supra

power they were willing of themselves;

8:4 Praying us with much intreaty that we would receive uirtutem uoluntarii fuerunt 8:4 cum multa exhortatione ob-

the gift, and take upon us the fellowship of the ministering to the saints.

secrantes nos gratiam et communicationem ministerii quod

8:5 And this they did, not as we hoped, but first gave their fit in sanctos 8:5 et non sicut sperauimus sed semet ipsos own selves to the Lord, and unto us by the will of God.

Epistola Pauli Ad Corinthios II

dederunt primum Domino deinde nobis per uoluntatem Dei

8:6 ita ut rogaremus Titum ut quemadmodum coepit ita et 8:6 Insomuch that we desired Titus, that as he had begun,

so he would also finish in you the same grace also.

perficiat in uos etiam gratiam istam 8:7 sed sicut in om- 8:7 Therefore, as ye abound in every thing, in faith, and utterance, and knowledge, and in all diligence, and in your love to us, see that ye abound in this grace also.

nibus abundatis fide et sermone et scientia et omni sollicitudine et caritate uestra in nos ut et in hac gratia abundetis 8:8 non quasi imperans dico sed per aliorum sollicitu- 8:8 I speak not by commandment, but by occasion of the

forwardness of others, and to prove the sincerity of your love.

dinem etiam uestrae caritatis ingenitum bonum conprobans 8:9 scitis enim gratiam Domini nostri Iesu Christi quoniam 8:9 For ye know the grace of our Lord Jesus Christ, that,

though he was rich, yet for your sakes he became poor, that ye through his poverty might be rich.

propter uos egenus factus est cum esset diues ut illius inopia

uos diuites essetis 8:10 et consilium in hoc do hoc enim 8:10 And herein I give my advice: for this is expedient for you, who have begun before, not only to do, but also to be forward a year ago.

uobis utile est qui non solum facere sed et uelle coepistis ab anno priore 8:11 nunc uero et facto perficite ut quemad- 8:11 Now therefore perform the doing of it; that as there

was a readiness to will, so there may be a performance also out of that which ye have.

modum promptus est animus uoluntatis ita sit et perficiendi

ex eo quod habetis 8:12 si enim uoluntas prompta est se- 8:12 For if there be first a willing mind, it is accepted according to that a man hath, and not according to that he hath not.

cundum id quod habet accepta est non secundum quod non habet 8:13 non enim ut aliis sit remissio uobis autem tri- 8:13 For I mean not that other men be eased, and ye bur-

dened:

bulatio sed ex aequalitate 8:14 in praesenti tempore uestra 8:14 But by an equality, that now at this time your abun-

dance may be a supply for their want, that their abundance also may be a supply for your want: that there may be

abundantia illorum inopiam suppleat ut et illorum abundan- equality: tia uestrae inopiae sit supplementum ut fiat aequalitas sicut

scriptum est 8:15 qui multum non abundauit et qui modi- 8:15 As it is written, He that had gathered much had noth-

ing over; and he that had gathered little had no lack.

cum non minorauit 8:16 gratias autem Deo qui dedit ean- 8:16 But thanks be to God, which put the same earnest

care into the heart of Titus for you.

dem sollicitudinem pro uobis in corde Titi 8:17 quoniam 8:17 For indeed he accepted the exhortation; but being

more forward, of his own accord he went unto you.

exhortationem quidem suscepit sed cum sollicitior esset sua uoluntate profectus est ad uos 8:18 misimus etiam cum illo 8:18 And we have sent with him the brother, whose praise

is in the gospel throughout all the churches;

fratrem cuius laus est in euangelio per omnes ecclesias 8:19 8:19 And not that only, but who was also chosen of the

churches to travel with us with this grace, which is administered by us to the glory of the same Lord, and declaration

non solum autem sed et ordinatus ab ecclesiis comes per- of your ready mind: egrinationis nostrae in hac gratia quae ministratur a nobis ad Domini gloriam et destinatam uoluntatem nostram 8:20 8:20 Avoiding this, that no man should blame us in this

abundance which is administered by us:

deuitantes hoc ne quis nos uituperet in hac plenitudine quae

ministratur a nobis 8:21 prouidemus enim bona non solum 8:21 Providing for honest things, not only in the sight of

the Lord, but also in the sight of men.

coram Deo sed etiam coram hominibus 8:22 misimus au- 8:22 And we have sent with them our brother, whom we

have oftentimes proved diligent in many things, but now much more diligent, upon the great confidence which I

tem cum illis et fratrem nostrum quem probauimus in multis have in you. saepe sollicitum esse nunc autem multo sollicitiorem confi-

Biblia Sacra Vulgata

8:23 Whether any do enquire of Titus, he is my partner and dentia multa in uos 8:23 siue pro Tito qui est socius meus et

fellowhelper concerning you: or our brethren be enquired of, they are the messengers of the churches, and the glory

of Christ. in uos adiutor siue fratres nostri apostoli ecclesiarum gloriae 8:24 Wherefore shew ye to them, and before the churches, Christi 8:24 ostensionem ergo quae est caritatis uestrae et

the proof of your love, and of our boasting on your behalf.

nostrae gloriae pro uobis in illos ostendite in faciem ecclesiarum

9:1 For as touching the ministering to the saints, it is su- 9:1 nam de ministerio quod fit in sanctos ex abundanti est perfluous for me to write to you: 9:2 For I know the forwardness of your mind, for which I mihi scribere uobis 9:2 scio enim promptum animum ueboast of you to them of Macedonia, that Achaia was ready a year ago; and your zeal hath provoked very many.

strum pro quo de uobis glorior apud Macedonas quoniam Achaia parata est ab anno praeterito et uestra aemulatio pro-

9:3 Yet have I sent the brethren, lest our boasting of you uocauit plurimos 9:3 misi autem fratres ut ne quod gloria-

should be in vain in this behalf; that, as I said, ye may be ready:

mur de uobis euacuetur in hac parte ut quemadmodum dixi

9:4 Lest haply if they of Macedonia come with me, and parati sitis 9:4 ne cum uenerint mecum Macedones et in-

find you unprepared, we (that we say not, ye) should be ashamed in this same confident boasting.

uenerint uos inparatos erubescamus nos ut non dicamus uos

9:5 Therefore I thought it necessary to exhort the brethren, in hac substantia 9:5 necessarium ergo existimaui rogare that they would go before unto you, and make up beforehand your bounty, whereof ye had notice before, that the same might be ready, as a matter of bounty, and not as of covetousness.

fratres ut praeueniant ad uos et praeparent repromissam benedictionem hanc paratam esse sic quasi benedictionem non

9:6 But this I say, He which soweth sparingly shall reap quasi auaritiam 9:6 hoc autem qui parce seminat parce et

also sparingly; and he which soweth bountifully shall reap also bountifully.

metet et qui seminat in benedictionibus de benedictionibus

9:7 Every man according as he purposeth in his heart, so et metet 9:7 unusquisque prout destinauit corde suo non ex

let him give; not grudgingly, or of necessity: for God loveth a cheerful giver.

tristitia aut ex necessitate hilarem enim datorem diligit Deus

9:8 And God is able to make all grace abound toward you; 9:8 potens est autem Deus omnem gratiam abundare facere in

that ye, always having all sufficiency in all things, may abound to every good work:

uobis ut in omnibus semper omnem sufficientiam habentes

9:9 (As it is written, He hath dispersed abroad; he hath abundetis in omne opus bonum 9:9 sicut scriptum est dis-

given to the poor: his righteousness remaineth for ever.

9:10 Now he that ministereth seed to the sower both min- persit dedit pauperibus iustitia eius manet in aeternum 9:10

ister bread for your food, and multiply your seed sown, and increase the fruits of your righteousness;)

qui autem administrat semen seminanti et panem ad manducandum praestabit et multiplicabit semen uestrum et auge-

9:11 Being enriched in every thing to all bountifulness, bit incrementa frugum iustitiae uestrae 9:11 ut in omnibus

which causeth through us thanksgiving to God.

locupletati abundetis in omnem simplicitatem quae opera-

9:12 For the administration of this service not only sup- tur per nos gratiarum actionem Deo 9:12 quoniam ministeplieth the want of the saints, but is abundant also by many thanksgivings unto God;

rium huius officii non solum supplet ea quae desunt sanctis sed etiam abundat per multas gratiarum actiones in Domino

9:13 Whiles by the experiment of this ministration they 9:13 per probationem ministerii huius glorificantes Deum in

glorify God for your professed subjection unto the gospel of Christ, and for your liberal distribution unto them, and

unto all men; oboedientia confessionis uestrae in euangelium Christi et

Epistola Pauli Ad Corinthios II

simplicitate communicationis in illos et in omnes 9:14 et 9:14 And by their prayer for you, which long after you for

the exceeding grace of God in you.

ipsorum obsecratione pro uobis desiderantium uos propter eminentem gratiam Dei in uobis 9:15 gratias Deo super in- 9:15 Thanks be unto God for his unspeakable gift. enarrabili dono eius

10:1 ipse autem ego Paulus obsecro uos per mansuetudi- 10:1 Now I Paul myself beseech you by the meekness and

gentleness of Christ, who in presence am base among you, but being absent am bold toward you:

nem et modestiam Christi qui in facie quidem humilis inter uos absens autem confido in uobis 10:2 rogo autem ne 10:2 But I beseech you, that I may not be bold when I am

present with that confidence, wherewith I think to be bold against some, which think of us as if we walked according

praesens audeam per eam confidentiam qua existimo audere to the flesh. in quosdam qui arbitrantur nos tamquam secundum carnem ambulemus 10:3 in carne enim ambulantes non secundum 10:3 For though we walk in the flesh, we do not war after

the flesh:

carnem militamus 10:4 nam arma militiae nostrae non car- 10:4 (For the weapons of our warfare are not carnal, but

mighty through God to the pulling down of strong holds;)

nalia sed potentia Deo ad destructionem munitionum consilia destruentes 10:5 et omnem altitudinem extollentem se 10:5 Casting down imaginations, and every high thing that

exalteth itself against the knowledge of God, and bringing into captivity every thought to the obedience of Christ;

aduersus scientiam Dei et in captiuitatem rediientes omnem intellectum in obsequium Christi 10:6 et in promptu haben- 10:6 And having in a readiness to revenge all disobedi-

ence, when your obedience is fulfilled.

tes ulcisci omnem inoboedientiam cum impleta fuerit uestra oboedientia 10:7 quae secundum faciem sunt uidete si quis 10:7 Do ye look on things after the outward appearance?

If any man trust to himself that he is Christ’s, let him of himself think this again, that, as he is Christ’s, even so are

confidit sibi Christi se esse hoc cogitet iterum apud se quia we Christ’s. sicut ipse Christi est ita et nos 10:8 nam et si amplius ali- 10:8 For though I should boast somewhat more of our au-

thority, which the Lord hath given us for edification, and not for your destruction, I should not be ashamed:

quid gloriatus fuero de potestate nostra quam dedit Dominus in aedificationem et non in destructionem uestram non erubescam 10:9 ut autem non existimer tamquam terrere uos 10:9 That I may not seem as if I would terrify you by let-

ters.

per epistulas 10:10 quoniam quidem epistulae inquiunt gra- 10:10 For his letters, say they, are weighty and power-

ful; but his bodily presence is weak, and his speech contemptible.

ues sunt et fortes praesentia autem corporis infirma et sermo contemptibilis 10:11 hoc cogitet qui eiusmodi est quia qua- 10:11 Let such an one think this, that, such as we are in

word by letters when we are absent, such will we be also in deed when we are present.

les sumus uerbo per epistulas absentes tales et praesentes in facto 10:12 non enim audemus inserere aut conparare 10:12 For we dare not make ourselves of the number, or

compare ourselves with some that commend themselves: but they measuring themselves by themselves, and com-

nos quibusdam qui se ipsos commendant sed ipsi in nobis paring themselves among themselves, are not wise. nosmet ipsos metientes et conparantes nosmet ipsos nobis

10:13 nos autem non in inmensum gloriabimur sed secun- 10:13 But we will not boast of things without our measure, but according to the measure of the rule which God hath distributed to us, a measure to reach even unto you.

dum mensuram regulae quam mensus est nobis Deus men-

suram pertingendi usque ad uos 10:14 non enim quasi non 10:14 For we stretch not ourselves beyond our measure, as though we reached not unto you: for we are come as far as to you also in preaching the gospel of Christ:

pertingentes ad uos superextendimus nos usque ad uos enim

Biblia Sacra Vulgata

10:15 Not boasting of things without our measure, that is, peruenimus in euangelio Christi 10:15 non in inmensum of other men’s labours; but having hope, when your faith is increased, that we shall be enlarged by you according to

our rule abundantly, gloriantes in alienis laboribus spem autem habentes cres-

centis fidei uestrae in uobis magnificari secundum regulam

10:16 To preach the gospel in the regions beyond you, and nostram in abundantiam 10:16 etiam in illa quae ultra uos

not to boast in another man’s line of things made ready to our hand.

sunt euangelizare non in aliena regula in his quae praeparata

10:17 But he that glorieth, let him glory in the Lord. sunt gloriari 10:17 qui autem gloriatur in Domino glorietur 10:18 For not he that commendeth himself is approved, 10:18 non enim qui se ipsum commendat ille probatus est sed

but whom the Lord commendeth.

quem Dominus commendat

11:1 Would to God ye could bear with me a little in my 11:1 utinam sustineretis modicum quid insipientiae meae

folly: and indeed bear with me. 11:2 For I am jealous over you with godly jealousy: for I sed et subportate me 11:2 aemulor enim uos Dei aemulahave espoused you to one husband, that I may present you as a chaste virgin to Christ.

tione despondi enim uos uni uiro uirginem castam exhibere

11:3 But I fear, lest by any means, as the serpent beguiled Christo 11:3 timeo autem ne sicut serpens Euam seduxit Eve through his subtilty, so your minds should be corrupted from the simplicity that is in Christ.

astutia sua ita corrumpantur sensus uestri et excidant a sim-

11:4 For if he that cometh preacheth another Jesus, whom plicitate quae est in Christo 11:4 nam si is qui uenit alium

we have not preached, or if ye receive another spirit, which ye have not received, or another gospel, which ye have not

accepted, ye might well bear with him. Christum praedicat quem non praedicauimus aut alium spiri-

tum accipitis quem non accepistis aut aliud euangelium quod

11:5 For I suppose I was not a whit behind the very non recepistis recte pateremini 11:5 existimo enim nihil me

chiefest apostles.

11:6 But though I be rude in speech, yet not in knowledge; minus fecisse magnis apostolis 11:6 et si inperitus sermone

but we have been throughly made manifest among you in all things.

sed non scientia in omnibus autem manifestatus sum uobis

11:7 Have I committed an offence in abasing myself that 11:7 aut numquid peccatum feci me ipsum humilians ut uos

ye might be exalted, because I have preached to you the gospel of God freely?

exaltemini quoniam gratis euangelium Dei euangelizaui uo-

11:8 I robbed other churches, taking wages of them, to do bis 11:8 alias ecclesias expoliaui accipiens stipendium ad

you service.

11:9 And when I was present with you, and wanted, I was ministerium uestrum 11:9 et cum essem apud uos et egerem

chargeable to no man: for that which was lacking to me the brethren which came from Macedonia supplied: and in all things I have kept myself from being burdensome unto you, and so will I keep myself.

nulli onerosus fui nam quod mihi deerat suppleuerunt fratres qui uenerunt a Macedonia et in omnibus sine onere me

11:10 As the truth of Christ is in me, no man shall stop me uobis seruaui et seruabo 11:10 est ueritas Christi in me quo-

of this boasting in the regions of Achaia.

niam haec gloria non infringetur in me in regionibus Achaiae

11:11 Wherefore? because I love you not? God knoweth. 11:11 quare quia non diligo uos Deus scit 11:12 quod au-

11:12 But what I do, that I will do, that I may cut off occasion from them which desire occasion; that wherein they glory, they may be found even as we.

tem facio et faciam ut amputem occasionem eorum qui uolunt occasionem ut in quo gloriantur inueniantur sicut et nos

11:13 For such are false apostles, deceitful workers, trans- 11:13 nam eiusmodi pseudoapostoli operarii subdoli transfiforming themselves into the apostles of Christ.

11:14 And no marvel; for Satan himself is transformed gurantes se in apostolos Christi 11:14 et non mirum ipse

into an angel of light. 11:15 Therefore it is no great thing if his ministers also be enim Satanas transfigurat se in angelum lucis 11:15 non est transformed as the ministers of righteousness; whose end shall be according to their works.

Epistola Pauli Ad Corinthios II

ergo magnum si ministri eius transfigurentur uelut ministri iustitiae quorum finis erit secundum opera ipsorum 11:16 11:16 I say again, Let no man think me a fool; if otherwise,

yet as a fool receive me, that I may boast myself a little.

iterum dico ne quis me putet insipientem alioquin uelut insi-

pientem accipite me ut et ego modicum quid glorier 11:17 11:17 That which I speak, I speak it not after the Lord, but

as it were foolishly, in this confidence of boasting.

quod loquor non loquor secundum Dominum sed quasi in insipientia in hac substantia gloriae 11:18 quoniam multi 11:18 Seeing that many glory after the flesh, I will glory

also.

gloriantur secundum carnem et ego gloriabor 11:19 liben- 11:19 For ye suffer fools gladly, seeing ye yourselves are

wise.

ter enim suffertis insipientes cum sitis ipsi sapientes 11:20 11:20 For ye suffer, if a man bring you into bondage, if

a man devour you, if a man take of you, if a man exalt himself, if a man smite you on the face.

sustinetis enim si quis uos in seruitutem rediiit si quis deuorat si quis accipit si quis extollitur si quis in faciem uos caedit 11:21 secundum ignobilitatem dico quasi nos infirmi 11:21 I speak as concerning reproach, as though we had

been weak. Howbeit whereinsoever any is bold, (I speak foolishly,) I am bold also.

fuerimus in quo quis audet in insipientia dico audeo et ego 11:22 Hebraei sunt et ego Israhelitae sunt et ego semen Ab- 11:22 Are they Hebrews? so am I. Are they Israelites? so

am I. Are they the seed of Abraham? so am I.

rahae sunt et ego 11:23 ministri Christi sunt minus sapiens 11:23 Are they ministers of Christ? (I speak as a fool) I am

more; in labours more abundant, in stripes above measure, in prisons more frequent, in deaths oft.

dico plus ego in laboribus plurimis in carceribus abundantius in plagis supra modum in mortibus frequenter 11:24 a 11:24 Of the Jews five times received I forty stripes save

one.

Iudaeis quinquies quadragenas una minus accepi 11:25 ter 11:25 Thrice was I beaten with rods, once was I stoned,

thrice I suffered shipwreck, a night and a day I have been in the deep;

uirgis caesus sum semel lapidatus sum ter naufragium feci nocte et die in profundo maris fui 11:26 in itineribus saepe 11:26 In journeyings often, in perils of waters, in perils

of robbers, in perils by mine own countrymen, in perils by the heathen, in perils in the city, in perils in the wilderness,

periculis fluminum periculis latronum periculis ex genere in perils in the sea, in perils among false brethren; periculis ex gentibus periculis in ciuitate periculis in solitudine periculis in mari periculis in falsis fratribus 11:27 in 11:27 In weariness and painfulness, in watchings often, in

hunger and thirst, in fastings often, in cold and nakedness.

labore et aerumna in uigiliis multis in fame et siti in ieiuniis multis in frigore et nuditate 11:28 praeter illa quae extrin- 11:28 Beside those things that are without, that which

cometh upon me daily, the care of all the churches.

secus sunt instantia mea cotidiana sollicitudo omnium ecclesiarum 11:29 quis infirmatur et non infirmor quis scandali- 11:29 Who is weak, and I am not weak? who is offended,

and I burn not?

zatur et ego non uror 11:30 si gloriari oportet quae infirmi- 11:30 If I must needs glory, I will glory of the things which

concern mine infirmities.

tatis meae sunt gloriabor 11:31 Deus et Pater Domini Iesu 11:31 The God and Father of our Lord Jesus Christ, which

is blessed for evermore, knoweth that I lie not.

scit qui est benedictus in saecula quod non mentior 11:32 11:32 In Damascus the governor under Aretas the king

kept the city of the Damascenes with a garrison, desirous to apprehend me:

Damasci praepositus gentis Aretae regis custodiebat ciuitatem Damascenorum ut me conprehenderet 11:33 et per fe- 11:33 And through a window in a basket was I let down

by the wall, and escaped his hands.

nestram in sporta dimissus sum per murum et effugi manus eius

Biblia Sacra Vulgata

12:1 It is not expedient for me doubtless to glory. I will 12:1 si gloriari oportet non expedit quidem ueniam aucome to visions and revelations of the Lord.

12:2 I knew a man in Christ above fourteen years ago, tem ad uisiones et reuelationes Domini 12:2 scio hominem

(whether in the body, I cannot tell; or whether out of the body, I cannot tell: God knoweth;) such an one caught up

to the third heaven. in Christo ante annos quattuordecim siue in corpore nescio

siue extra corpus nescio Deus scit raptum eiusmodi usque

12:3 And I knew such a man, (whether in the body, or out ad tertium caelum 12:3 et scio huiusmodi hominem siue in

of the body, I cannot tell: God knoweth;)

12:4 How that he was caught up into paradise, and heard corpore siue extra corpus nescio Deus scit 12:4 quoniam

unspeakable words, which it is not lawful for a man to utter.

raptus est in paradisum et audiuit arcana uerba quae non li-

12:5 Of such an one will I glory: yet of myself I will not cet homini loqui 12:5 pro eiusmodi gloriabor pro me autem

glory, but in mine infirmities.

12:6 For though I would desire to glory, I shall not be a nihil gloriabor nisi in infirmitatibus meis 12:6 nam et si uo-

fool; for I will say the truth: but now I forbear, lest any man should think of me above that which he seeth me to

be, or that he heareth of me. luero gloriari non ero insipiens ueritatem enim dicam parco

autem ne quis in me existimet supra id quod uidet me aut

12:7 And lest I should be exalted above measure through audit ex me 12:7 et ne magnitudo reuelationum extollat me

the abundance of the revelations, there was given to me a thorn in the flesh, the messenger of Satan to buffet me, lest

I should be exalted above measure. datus est mihi stimulus carnis meae angelus Satanae ut me

12:8 For this thing I besought the Lord thrice, that it might colaphizet 12:8 propter quod ter Dominum rogaui ut disdepart from me.

12:9 And he said unto me, My grace is sufficient for thee: cederet a me 12:9 et dixit mihi sufficit tibi gratia mea nam

for my strength is made perfect in weakness. Most gladly therefore will I rather glory in my infirmities, that the

power of Christ may rest upon me. uirtus in infirmitate perficitur libenter igitur gloriabor in in-12:10 Therefore I take pleasure in infirmities, in re- firmitatibus meis ut inhabitet in me uirtus Christi 12:10 pro-

proaches, in necessities, in persecutions, in distresses for Christ’s sake: for when I am weak, then am I strong.

pter quod placeo mihi in infirmitatibus in contumeliis in necessitatibus in persecutionibus in angustiis pro Christo cum

12:11 I am become a fool in glorying; ye have compelled enim infirmor tunc potens sum 12:11 factus sum insipiens

me: for I ought to have been commended of you: for in nothing am I behind the very chiefest apostles, though I

be nothing. uos me coegistis ego enim debui a uobis commendari nihil

enim minus fui ab his qui sunt supra modum apostoli tametsi

12:12 Truly the signs of an apostle were wrought among nihil sum 12:12 signa tamen apostoli facta sunt super uos in

you in all patience, in signs, and wonders, and mighty deeds.

omni patientia signis et prodiiiis et uirtutibus 12:13 quid est

12:13 For what is it wherein ye were inferior to other churches, except it be that I myself was not burdensome to you? forgive me this wrong.

enim quod minus habuistis prae ceteris ecclesiis nisi quod

12:14 Behold, the third time I am ready to come to you; ego ipse non grauaui uos donate mihi hanc iniuriam 12:14

and I will not be burdensome to you: for I seek not your’s but you: for the children ought not to lay up for the par-

ents, but the parents for the children. ecce tertio hoc paratus sum uenire ad uos et non ero grauis

uobis non enim quaero quae uestra sunt sed uos nec enim

12:15 And I will very gladly spend and be spent for you; debent filii parentibus thesaurizare sed parentes filiis 12:15

though the more abundantly I love you, the less I be loved.

ego autem libentissime inpendam et superinpendar ipse pro

12:16 But be it so, I did not burden you: nevertheless, animabus uestris licet plus uos diligens minus diligar 12:16

being crafty, I caught you with guile.

sed esto ego uos non grauaui sed cum essem astutus dolo uos

12:17 Did I make a gain of you by any of them whom I cepi 12:17 numquid per aliquem eorum quos misi ad uos

sent unto you?

Epistola Pauli Ad Corinthios II

circumueni uos 12:18 rogaui Titum et misi cum illo fratrem 12:18 I desired Titus, and with him I sent a brother. Did

Titus make a gain of you? walked we not in the same spirit? walked we not in the same steps?

numquid Titus uos circumuenit nonne eodem spiritu ambulauimus nonne hisdem uestigiis 12:19 olim putatis quod ex- 12:19 Again, think ye that we excuse ourselves unto you?

we speak before God in Christ: but we do all things, dearly beloved, for your edifying.

cusemus nos apud uos coram Deo in Christo loquimur omnia autem carissimi propter uestram aedificationem 12:20 12:20 For I fear, lest, when I come, I shall not find you

such as I would, and that I shall be found unto you such as ye would not: lest there be debates, envyings, wraths,

timeo enim ne forte cum uenero non quales uolo inueniam strifes, backbitings, whisperings, swellings, tumults:

uos et ego inueniar a uobis qualem non uultis ne forte contentiones aemulationes animositates dissensiones detractiones susurrationes inflationes seditiones sint inter uos 12:21 12:21 And lest, when I come again, my God will humble

me among you, and that I shall bewail many which have sinned already, and have not repented of the uncleanness and fornication and lasciviousness which they have committed.

ne iterum cum uenero humiliet me Deus apud uos et lugeam multos ex his qui ante peccauerunt et non egerunt paenitentiam super inmunditia et fornicatione et inpudicitia quam gesserunt

13:1 ecce tertio hoc uenio ad uos in ore duorum uel trium 13:1 This is the third time I am coming to you. In the

mouth of two or three witnesses shall every word be established.

testium stabit omne uerbum 13:2 praedixi et praedico ut

13:2 I told you before, and foretell you, as if I were present, the second time; and being absent now I write to them which heretofore have sinned, and to all other, that, if I come again, I will not spare:

praesens bis et nunc absens his qui ante peccauerunt et cete-

ris omnibus quoniam si uenero iterum non parcam 13:3 an 13:3 Since ye seek a proof of Christ speaking in me, which

to you-ward is not weak, but is mighty in you.

experimentum quaeritis eius qui in me loquitur Christi qui in uos non infirmatur sed potens est in uobis 13:4 nam et si 13:4 For though he was crucified through weakness, yet he

liveth by the power of God. For we also are weak in him, but we shall live with him by the power of God toward

crucifixus est ex infirmitate sed uiuit ex uirtute Dei nam et you. nos infirmi sumus in illo sed uiuemus cum eo ex uirtute Dei in uobis 13:5 uosmet ipsos temptate si estis in fide ipsi uos 13:5 Examine yourselves, whether ye be in the faith; prove

your own selves. Know ye not your own selves, how that Jesus Christ is in you, except ye be reprobates?

probate an non cognoscitis uos ipsos quia Christus Iesus in uobis est nisi forte reprobi estis 13:6 spero autem quod co- 13:6 But I trust that ye shall know that we are not repro-

bates.

gnoscetis quia nos non sumus reprobi 13:7 oramus autem 13:7 Now I pray to God that ye do no evil; not that we

should appear approved, but that ye should do that which is honest, though we be as reprobates.

Deum ut nihil mali faciatis non ut nos probati pareamus sed ut uos quod bonum est faciatis nos autem ut reprobi simus 13:8 non enim possumus aliquid aduersus ueritatem sed pro 13:8 For we can do nothing against the truth, but for the

truth.

ueritate 13:9 gaudemus enim quando nos infirmi sumus uos 13:9 For we are glad, when we are weak, and ye are strong:

and this also we wish, even your perfection.

autem potentes estis hoc et oramus uestram consummatio-

nem 13:10 ideo haec absens scribo ut non praesens durius 13:10 Therefore I write these things being absent, lest being present I should use sharpness, according to the power which the Lord hath given me to edification, and not to

agam secundum potestatem quam Dominus dedit mihi in ae- destruction. dificationem et non in destructionem 13:11 de cetero frat- 13:11 Finally, brethren, farewell. Be perfect, be of good

comfort, be of one mind, live in peace; and the God of love and peace shall be with you.

Biblia Sacra Vulgata

res gaudete perfecti estote exhortamini idem sapite pacem

13:12 Greet one another with an holy kiss. habete et Deus dilectionis et pacis erit uobiscum 13:12 sa-

lutate inuicem in osculo sancto salutant uos sancti omnes

13:13 All the saints salute you. 13:13 gratia Domini nostri Iesu Christi et caritas Dei et communicatio Sancti Spiritus cum omnibus uobis amen

Epistola Pauli Ad Galatas

1:1 Paulus apostolus non ab hominibus neque per homi- 1:1 Paul, an apostle, (not of men, neither by man, but by Jesus Christ, and God the Father, who raised him from the dead;)

nem sed per Iesum Christum et Deum Patrem qui suscitauit eum a mortuis 1:2 et qui mecum sunt omnes fratres eccle- 1:2 And all the brethren which are with me, unto the

churches of Galatia:

siis Galatiae 1:3 gratia uobis et pax a Deo Patre et Domino 1:3 Grace be to you and peace from God the Father, and

from our Lord Jesus Christ,

nostro Iesu Christo 1:4 qui dedit semet ipsum pro pecca- 1:4 Who gave himself for our sins, that he might deliver us from this present evil world, according to the will of God and our Father:

tis nostris ut eriperet nos de praesenti saeculo nequam secundum uoluntatem Dei et Patris nostri 1:5 cui est gloria 1:5 To whom be glory for ever and ever. Amen. in saecula saeculorum amen 1:6 miror quod sic tam cito 1:6 I marvel that ye are so soon removed from him that

called you into the grace of Christ unto another gospel:

transferimini ab eo qui uos uocauit in gratiam Christi in aliud euangelium 1:7 quod non est aliud nisi sunt aliqui qui 1:7 Which is not another; but there be some that trouble

you, and would pervert the gospel of Christ.

uos conturbant et uolunt conuertere euangelium Christi 1:8 1:8 But though we, or an angel from heaven, preach any

other gospel unto you than that which we have preached unto you, let him be accursed.

sed licet nos aut angelus de caelo euangelizet uobis praeterquam quod euangelizauimus uobis anathema sit 1:9 sicut 1:9 As we said before, so say I now again, if any man

preach any other gospel unto you than that ye have received, let him be accursed.

praediximus et nunc iterum dico si quis uobis euangelizauerit praeter id quod accepistis anathema sit 1:10 modo enim 1:10 For do I now persuade men, or God? or do I seek to

please men? for if I yet pleased men, I should not be the servant of Christ.

hominibus suadeo aut Deo aut quaero hominibus placere si

adhuc hominibus placerem Christi seruus non essem 1:11 1:11 But I certify you, brethren, that the gospel which was

preached of me is not after man.

notum enim uobis facio fratres euangelium quod euangeliza-

tum est a me quia non est secundum hominem 1:12 neque 1:12 For I neither received it of man, neither was I taught

it, but by the revelation of Jesus Christ.

enim ego ab homine accepi illud neque didici sed per reuelationem Iesu Christi 1:13 audistis enim conuersationem 1:13 For ye have heard of my conversation in time past in

the Jews’ religion, how that beyond measure I persecuted the church of God, and wasted it:

meam aliquando in iudaismo quoniam supra modum persequebar ecclesiam Dei et expugnabam illam 1:14 et profi- 1:14 And profited in the Jews’ religion above many my

equals in mine own nation, being more exceedingly zealous of the traditions of my fathers.

ciebam in iudaismo supra multos coetaneos in genere meo abundantius aemulator existens paternarum mearum traditionum 1:15 cum autem placuit ei qui me segregauit de 1:15 But when it pleased God, who separated me from my

mother’s womb, and called me by his grace,

utero matris meae et uocauit per gratiam suam 1:16 ut re- 1:16 To reveal his Son in me, that I might preach him

among the heathen; immediately I conferred not with flesh and blood:

uelaret Filium suum in me ut euangelizarem illum in gentibus continuo non adquieui carni et sanguini 1:17 neque 1:17 Neither went I up to Jerusalem to them which were

apostles before me; but I went into Arabia, and returned again unto Damascus.

ueni Hierosolyma ad antecessores meos apostolos sed abii in Arabiam et iterum reuersus sum Damascum 1:18 deinde 1:18 Then after three years I went up to Jerusalem to see

Peter, and abode with him fifteen days.

post annos tres ueni Hierosolyma uidere Petrum et mansi apud eum diebus quindecim 1:19 alium autem apostolorum 1:19 But other of the apostles saw I none, save James the

Lord’s brother.

Biblia Sacra Vulgata

1:20 Now the things which I write unto you, behold, be- uidi neminem nisi Iacobum fratrem Domini 1:20 quae au-

fore God, I lie not. 1:21 Afterwards I came into the regions of Syria and Cili- tem scribo uobis ecce coram Deo quia non mentior 1:21 cia;

1:22 And was unknown by face unto the churches of Ju- deinde ueni in partes Syriae et Ciliciae 1:22 eram autem

daea which were in Christ:

1:23 But they had heard only, That he which persecuted ignotus facie ecclesiis Iudaeae quae erant in Christo 1:23

us in times past now preacheth the faith which once he destroyed.

tantum autem auditum habebant quoniam qui persequebatur nos aliquando nunc euangelizat fidem quam aliquando ex-

1:24 And they glorified God in me. pugnabat 1:24 et in me clarificabant Deum

2:1 Then fourteen years after I went up again to Jerusalem 2:1 deinde post annos quattuordecim iterum ascendi Hiewith Barnabas, and took Titus with me also. 2:2 And I went up by revelation, and communicated unto rosolyma cum Barnaba adsumpto et Tito 2:2 ascendi authem that gospel which I preach among the Gentiles, but privately to them which were of reputation, lest by any

means I should run, or had run, in vain. tem secundum reuelationem et contuli cum illis euangelium

quod praedico in gentibus seorsum autem his qui uideban-

2:3 But neither Titus, who was with me, being a Greek, tur ne forte in uacuum currerem aut cucurrissem 2:3 sed

was compelled to be circumcised:

neque Titus qui mecum erat cum esset gentilis conpulsus

2:4 And that because of false brethren unawares brought est circumcidi 2:4 sed propter subintroductos falsos fratres

in, who came in privily to spy out our liberty which we have in Christ Jesus, that they might bring us into bondage:

qui subintroierunt explorare libertatem nostram quam habe-

2:5 To whom we gave place by subjection, no, not for an mus in Christo Iesu ut nos in seruitutem rediierent 2:5 qui-

hour; that the truth of the gospel might continue with you.

bus neque ad horam cessimus subiectioni ut ueritas euan-

2:6 But of these who seemed to be somewhat, (whatsoever gelii permaneat apud uos 2:6 ab his autem qui uidebantur

they were, it maketh no matter to me: God accepteth no man’s person:) for they who seemed to be somewhat in

conference added nothing to me: esse aliquid quales aliquando fuerint nihil mea interest De-

us personam hominis non accipit mihi enim qui uidebantur

2:7 But contrariwise, when they saw that the gospel of the nihil contulerunt 2:7 sed e contra cum uidissent quod cre-

uncircumcision was committed unto me, as the gospel of the circumcision was unto Peter;

ditum est mihi euangelium praeputii sicut Petro circumci-

2:8 (For he that wrought effectually in Peter to the apos- sionis 2:8 qui enim operatus est Petro in apostolatum cir-

tleship of the circumcision, the same was mighty in me toward the Gentiles:)

cumcisionis operatus est et mihi inter gentes 2:9 et cum

2:9 And when James, Cephas, and John, who seemed to be pillars, perceived the grace that was given unto me, they gave to me and Barnabas the right hands of fellowship; that we should go unto the heathen, and they unto the circumcision.

cognouissent gratiam quae data est mihi Iacobus et Cephas et Iohannes qui uidebantur columnae esse dextras dederunt mihi et Barnabae societatis ut nos in gentes ipsi autem in

2:10 Only they would that we should remember the poor; circumcisionem 2:10 tantum ut pauperum memores esse-

the same which I also was forward to do.

2:11 But when Peter was come to Antioch, I withstood mus quod etiam sollicitus fui hoc ipsum facere 2:11 cum

him to the face, because he was to be blamed.

autem uenisset Cephas Antiochiam in faciem ei restiti quia

2:12 For before that certain came from James, he did eat reprehensibilis erat 2:12 prius enim quam uenirent quidam

with the Gentiles: but when they were come, he withdrew and separated himself, fearing them which were of the cir-

cumcision. ab Iacobo cum gentibus edebat cum autem uenissent subtrahebat et segregabat se timens eos qui ex circumcisione erant

Epistola Pauli Ad Galatas

2:13 et simulationi eius consenserunt ceteri Iudaei ita ut et 2:13 And the other Jews dissembled likewise with him;

insomuch that Barnabas also was carried away with their dissimulation.

Barnabas duceretur ab eis in illa simulatione 2:14 sed cum

2:14 But when I saw that they walked not uprightly according to the truth of the gospel, I said unto Peter before them all, If thou, being a Jew, livest after the manner of Gentiles, and not as do the Jews, why compellest thou the Gentiles to live as do the Jews?

uidissem quod non recte ambularent ad ueritatem euangelii dixi Cephae coram omnibus si tu cum Iudaeus sis gentiliter et non iudaice uiuis quomodo gentes cogis iudaizare 2:15 2:15 We who are Jews by nature, and not sinners of the

Gentiles,

nos natura Iudaei et non ex gentibus peccatores 2:16 scien- 2:16 Knowing that a man is not justified by the works of

the law, but by the faith of Jesus Christ, even we have believed in Jesus Christ, that we might be justified by the faith of Christ, and not by the works of the law: for by the works of the law shall no flesh be justified.

tes autem quod non iustificatur homo ex operibus legis nisi per fidem Iesu Christi et nos in Christo Iesu credidimus ut iustificemur ex fide Christi et non ex operibus legis propter quod ex operibus legis non iustificabitur omnis caro 2:17 2:17 But if, while we seek to be justified by Christ, we

ourselves also are found sinners, is therefore Christ the minister of sin? God forbid.

quod si quaerentes iustificari in Christo inuenti sumus et ipsi peccatores numquid Christus peccati minister est absit 2:18 2:18 For if I build again the things which I destroyed, I

make myself a transgressor.

si enim quae destruxi haec iterum aedifico praeuaricatorem me constituo 2:19 ego enim per legem legi mortuus sum ut 2:19 For I through the law am dead to the law, that I might

live unto God.

Deo uiuam Christo confixus sum cruci 2:20 uiuo autem iam 2:20 I am crucified with Christ: neverthless I live; yet not I, but Christ liveth in me: and the life which I now live in the flesh I live by the faith of the Son of God, who loved

non ego uiuit uero in me Christus quod autem nunc uiuo in me, and gave himself for me. carne in fide uiuo Filii Dei qui dilexit me et tradidit se ip-

sum pro me 2:21 non abicio gratiam Dei si enim per legem 2:21 I do not frustrate the grace of God: for if righteous-

ness come by the law, then Christ is dead in vain.

iustitia ergo Christus gratis mortuus est

3:1 o insensati Galatae quis uos fascinauit ante quorum 3:1 O foolish Galatians, who hath bewitched you, that ye

should not obey the truth, before whose eyes Jesus Christ hath been evidently set forth, crucified among you?

oculos Iesus Christus proscriptus est crucifixus 3:2 hoc so-

3:2 This only would I learn of you, Received ye the Spirit by the works of the law, or by the hearing of faith?

lum uolo a uobis discere ex operibus legis Spiritum accepistis an ex auditu fidei 3:3 sic stulti estis cum Spiritu coe- 3:3 Are ye so foolish? having begun in the Spirit, are ye

now made perfect by the flesh?

peritis nunc carne consummamini 3:4 tanta passi estis sine 3:4 Have ye suffered so many things in vain? if it be yet

in vain.

causa si tamen sine causa 3:5 qui ergo tribuit uobis Spi- 3:5 He therefore that ministereth to you the Spirit, and

worketh miracles among you, doeth he it by the works of the law, or by the hearing of faith?

ritum et operatur uirtutes in uobis ex operibus legis an ex auditu fidei 3:6 sicut Abraham credidit Deo et reputatum 3:6 Even as Abraham believed God, and it was accounted

to him for righteousness.

est ei ad iustitiam 3:7 cognoscitis ergo quia qui ex fide sunt 3:7 Know ye therefore that they which are of faith, the

same are the children of Abraham.

hii sunt filii Abrahae 3:8 prouidens autem scriptura quia ex 3:8 And the scripture, foreseeing that God would justify

the heathen through faith, preached before the gospel unto Abraham, saying, In thee shall all nations be blessed.

fide iustificat gentes Deus praenuntiauit Abrahae quia be-

nedicentur in te omnes gentes 3:9 igitur qui ex fide sunt 3:9 So then they which be of faith are blessed with faithful

Abraham.

benedicentur cum fideli Abraham 3:10 quicumque enim ex 3:10 For as many as are of the works of the law are under

the curse: for it is written, Cursed is every one that continueth not in all things which are written in the book of the

operibus legis sunt sub maledicto sunt scriptum est enim ma- law to do them.

Biblia Sacra Vulgata

ledictus omnis qui non permanserit in omnibus quae scripta

3:11 But that no man is justified by the law in the sight of sunt in libro legis ut faciat ea 3:11 quoniam autem in lege

God, it is evident: for, The just shall live by faith.

nemo iustificatur apud Deum manifestum est quia iustus ex

3:12 And the law is not of faith: but, The man that doeth fide uiuit 3:12 lex autem non est ex fide sed qui fecerit ea

them shall live in them.

3:13 Christ hath redeemed us from the curse of the law, uiuet in illis 3:13 Christus nos redeeit de maledicto legis

being made a curse for us: for it is written, Cursed is every one that hangeth on a tree:

factus pro nobis maledictum quia scriptum est maledictus

3:14 That the blessing of Abraham might come on the omnis qui pendet in ligno 3:14 ut in gentibus benedictio

Gentiles through Jesus Christ; that we might receive the promise of the Spirit through faith.

Abrahae fieret in Christo Iesu ut pollicitationem Spiritus ac-

3:15 Brethren, I speak after the manner of men; Though cipiamus per fidem 3:15 fratres secundum hominem dico it be but a man’s covenant, yet if it be confirmed, no man disannulleth, or addeth thereto.

tamen hominis confirmatum testamentum nemo spernit aut

3:16 Now to Abraham and his seed were the promises superordinat 3:16 Abrahae dictae sunt promissiones et se-

made. He saith not, And to seeds, as of many; but as of one, And to thy seed, which is Christ.

mini eius non dicit et seminibus quasi in multis sed quasi

3:17 And this I say, that the covenant, that was confirmed in uno et semini tuo qui est Christus 3:17 hoc autem dico before of God in Christ, the law, which was four hundred and thirty years after, cannot disannul, that it should make

the promise of none effect. testamentum confirmatum a Deo quae post quadringentos et

triginta annos facta est lex non irritam facit ad euacuandam

3:18 For if the inheritance be of the law, it is no more of promissionem 3:18 nam si ex lege hereditas iam non ex re-

promise: but God gave it to Abraham by promise.

promissione Abrahae autem per promissionem donauit Deus

3:19 Wherefore then serveth the law? It was added be- 3:19 quid igitur lex propter transgressiones posita est donec

cause of transgressions, till the seed should come to whom the promise was made; and it was ordained by angels in

the hand of a mediator. ueniret semen cui promiserat ordinata per angelos in manu 3:20 Now a mediator is not a mediator of one, but God is mediatoris 3:20 mediator autem unius non est Deus autem

one.

3:21 Is the law then against the promises of God? God unus est 3:21 lex ergo aduersus promissa Dei absit si enim

forbid: for if there had been a law given which could have given life, verily righteousness should have been by the

law. data esset lex quae posset uiuificare uere ex lege esset iustitia 3:22 But the scripture hath concluded all under sin, that 3:22 sed conclusit scriptura omnia sub peccato ut promis-

the promise by faith of Jesus Christ might be given to them that believe.

sio ex fide Iesu Christi daretur credentibus 3:23 prius autem

3:23 But before faith came, we were kept under the law, shut up unto the faith which should afterwards be revealed.

quam ueniret fides sub lege custodiebamur conclusi in eam

3:24 Wherefore the law was our schoolmaster to bring us fidem quae reuelanda erat 3:24 itaque lex pedagogus noster

unto Christ, that we might be justified by faith.

3:25 But after that faith is come, we are no longer under a fuit in Christo ut ex fide iustificemur 3:25 at ubi uenit fides

schoolmaster. 3:26 For ye are all the children of God by faith in Christ iam non sumus sub pedagogo 3:26 omnes enim filii Dei es-Jesus. 3:27 For as many of you as have been baptized into Christ tis per fidem in Christo Iesu 3:27 quicumque enim in Chhave put on Christ.

3:28 There is neither Jew nor Greek, there is neither bond risto baptizati estis Christum induistis 3:28 non est Iudaeus

nor free, there is neither male nor female: for ye are all one in Christ Jesus.

neque Graecus non est seruus neque liber non est masculus neque femina omnes enim uos unum estis in Christo Iesu

3:29 And if ye be Christ’s, then are ye Abraham’s seed, 3:29 si autem uos Christi ergo Abrahae semen estis secun-

and heirs according to the promise.

Epistola Pauli Ad Galatas

dum promissionem heredes

4:1 dico autem quanto tempore heres paruulus est nihil 4:1 Now I say, That the heir, as long as he is a child, dif-

fereth nothing from a servant, though he be lord of all;

differt seruo cum sit dominus omnium 4:2 sed sub tutori- 4:2 But is under tutors and governors until the time ap-

pointed of the father.

bus est et actoribus usque ad praefinitum tempus a patre 4:3 4:3 Even so we, when we were children, were in bondage

under the elements of the world:

ita et nos cum essemus paruuli sub elementis mundi eramus seruientes 4:4 at ubi uenit plenitudo temporis misit Deus 4:4 But when the fulness of the time was come, God sent

forth his Son, made of a woman, made under the law,

Filium suum factum ex muliere factum sub lege 4:5 ut eos 4:5 To redeem them that were under the law, that we might

receive the adoption of sons.

qui sub lege erant rediieret ut adoptionem filiorum reciperemus 4:6 quoniam autem estis filii misit Deus Spiritum Filii 4:6 And because ye are sons, God hath sent forth the Spirit

of his Son into your hearts, crying, Abba, Father.

sui in corda nostra clamantem Abba Pater 4:7 itaque iam 4:7 Wherefore thou art no more a servant, but a son; and

if a son, then an heir of God through Christ.

non es seruus sed filius quod si filius et heres per Deum 4:8 4:8 Howbeit then, when ye knew not God, ye did service

unto them which by nature are no gods.

sed tunc quidem ignorantes Deum his qui natura non sunt dii seruiebatis 4:9 nunc autem cum cognoueritis Deum immo 4:9 But now, after that ye have known God, or rather are

known of God, how turn ye again to the weak and beggarly elements, whereunto ye desire again to be in bondage?

cogniti sitis a Deo quomodo conuertimini iterum ad infirma et egena elementa quibus denuo seruire uultis 4:10 dies ob- 4:10 Ye observe days, and months, and times, and years. seruatis et menses et tempora et annos 4:11 timeo uos ne 4:11 I am afraid of you, lest I have bestowed upon you

labour in vain.

forte sine causa laborauerim in uobis 4:12 estote sicut et 4:12 Brethren, I beseech you, be as I am; for I am as ye

are: ye have not injured me at all.

ego quia et ego sicut uos fratres obsecro uos nihil me laesistis 4:13 scitis autem quia per infirmitatem carnis euan- 4:13 Ye know how through infirmity of the flesh I

preached the gospel unto you at the first.

gelizaui uobis iam pridem 4:14 et temptationem uestram in 4:14 And my temptation which was in my flesh ye de-

spised not, nor rejected; but received me as an angel of God, even as Christ Jesus.

carne mea non spreuistis neque respuistis sed sicut angelum Dei excepistis me sicut Christum Iesum 4:15 ubi est ergo 4:15 Where is then the blessedness ye spake of? for I bear

you record, that, if it had been possible, ye would have plucked out your own eyes, and have given them to me.

beatitudo uestra testimonium enim perhibeo uobis quia si fieri posset oculos uestros eruissetis et dedissetis mihi 4:16 4:16 Am I therefore become your enemy, because I tell

you the truth?

ergo inimicus uobis factus sum uerum dicens uobis 4:17 ae- 4:17 They zealously affect you, but not well; yea, they

would exclude you, that ye might affect them.

mulantur uos non bene sed excludere uos uolunt ut illos aemulemini 4:18 bonum autem aemulamini in bono semper 4:18 But it is good to be zealously affected always in a

good thing, and not only when I am present with you.

et non tantum cum praesens sum apud uos 4:19 filioli mei 4:19 My little children, of whom I travail in birth again

until Christ be formed in you,

quos iterum parturio donec formetur Christus in uobis 4:20 4:20 I desire to be present with you now, and to change

my voice; for I stand in doubt of you.

uellem autem esse apud uos modo et mutare uocem meam quoniam confundor in uobis 4:21 dicite mihi qui sub lege 4:21 Tell me, ye that desire to be under the law, do ye not

hear the law?

uultis esse legem non legistis 4:22 scriptum est enim quo- 4:22 For it is written, that Abraham had two sons, the one

by a bondmaid, the other by a freewoman.

niam Abraham duos filios habuit unum de ancilla et unum

Biblia Sacra Vulgata

4:23 But he who was of the bondwoman was born after de libera 4:23 sed qui de ancilla secundum carnem natus

the flesh; but he of the freewoman was by promise.

4:24 Which things are an allegory: for these are the two est qui autem de libera per repromissionem 4:24 quae sunt

covenants; the one from the mount Sinai, which gendereth to bondage, which is Agar.

per allegoriam dicta haec enim sunt duo testamenta unum quidem a monte Sina in seruitutem generans quae est Agar

4:25 For this Agar is mount Sinai in Arabia, and answereth 4:25 Sina enim mons est in Arabia qui coniunctus est ei quae

to Jerusalem which now is, and is in bondage with her children.

nunc est Hierusalem et seruit cum filiis eius 4:26 illa autem

4:26 But Jerusalem which is above is free, which is the mother of us all.

quae sursum est Hierusalem libera est quae est mater nostra

4:27 For it is written, Rejoice, thou barren that bearest not; 4:27 scriptum est enim laetare sterilis quae non paris erumpe break forth and cry, thou that travailest not: for the desolate hath many more children than she which hath an hus-

band. et exclama quae non parturis quia multi filii desertae magis 4:28 Now we, brethren, as Isaac was, are the children of quam eius quae habet uirum 4:28 nos autem fratres secun-

promise. 4:29 But as then he that was born after the flesh persecuted dum Isaac promissionis filii sumus 4:29 sed quomodo tunc him that was born after the Spirit, even so it is now.

qui secundum carnem natus fuerat persequebatur eum qui

4:30 Nevertheless what saith the scripture? Cast out the secundum spiritum ita et nunc 4:30 sed quid dicit scriptura

bondwoman and her son: for the son of the bondwoman shall not be heir with the son of the freewoman.

eice ancillam et filium eius non enim heres erit filius ancillae

4:31 So then, brethren, we are not children of the bond- cum filio liberae 4:31 itaque fratres non sumus ancillae filii

woman, but of the free.

sed liberae qua libertate nos Christus liberauit

5:1 Stand fast therefore in the liberty wherewith Christ 5:1 state et nolite iterum iugo seruitutis contineri 5:2 ecce

hath made us free, and be not entangled again with the yoke of bondage. 5:2 Behold, I Paul say unto you, that if ye be circumcised, Christ shall profit you nothing.

ego Paulus dico uobis quoniam si circumcidamini Chris-

5:3 For I testify again to every man that is circumcised, tus uobis nihil prodeeit 5:3 testificor autem rursum omni

that he is a debtor to do the whole law.

homini circumcidenti se quoniam debitor est uniuersae le-

5:4 Christ is become of no effect unto you, whosoever of gis faciendae 5:4 euacuati estis a Christo qui in lege iu-

you are justified by the law; ye are fallen from grace.

5:5 For we through the Spirit wait for the hope of righ- stificamini a gratia excidistis 5:5 nos enim spiritu ex fide

teousness by faith. 5:6 For in Jesus Christ neither circumcision availeth any spem iustitiae expectamus 5:6 nam in Christo Iesu neque thing, nor uncircumcision; but faith which worketh by love.

circumcisio aliquid ualet neque praeputium sed fides quae

5:7 Ye did run well; who did hinder you that ye should not per caritatem operatur 5:7 currebatis bene quis uos inpe-

obey the truth?

5:8 This persuasion cometh not of him that calleth you. diuit ueritati non oboedire 5:8 persuasio non est ex eo qui 5:9 A little leaven leaveneth the whole lump. uocat uos 5:9 modicum fermentum totam massam corrum-5:10 I have confidence in you through the Lord, that ye pit 5:10 ego confido in uobis in Domino quod nihil aliud

will be none otherwise minded: but he that troubleth you shall bear his judgment, whosoever he be.

sapietis qui autem conturbat uos portabit iudicium quicum-

5:11 And I, brethren, if I yet preach circumcision, why do que est ille 5:11 ego autem fratres si circumcisionem ad-I yet suffer persecution? then is the offence of the cross ceased.

huc praedico quid adhuc persecutionem patior ergo euacua-

5:12 I would they were even cut off which trouble you. tum est scandalum crucis 5:12 utinam et abscidantur qui 5:13 For, brethren, ye have been called unto liberty; only uos conturbant 5:13 uos enim in libertatem uocati estis fra-

use not liberty for an occasion to the flesh, but by love serve one another.

Epistola Pauli Ad Galatas

tres tantum ne libertatem in occasionem detis carnis sed per caritatem seruite inuicem 5:14 omnis enim lex in uno ser- 5:14 For all the law is fulfilled in one word, even in this;

Thou shalt love thy neighbour as thyself.

mone impletur diliges proximum tuum sicut te ipsum 5:15 5:15 But if ye bite and devour one another, take heed that

ye be not consumed one of another.

quod si inuicem mordetis et comeditis uidete ne ab inuicem consumamini 5:16 dico autem spiritu ambulate et deside- 5:16 This I say then, Walk in the Spirit, and ye shall not

fulfil the lust of the flesh.

rium carnis non perficietis 5:17 caro enim concupiscit ad- 5:17 For the flesh lusteth against the Spirit, and the Spirit against the flesh: and these are contrary the one to the other: so that ye cannot do the things that ye would.

uersus spiritum spiritus autem aduersus carnem haec enim inuicem aduersantur ut non quaecumque uultis illa faciatis 5:18 quod si spiritu ducimini non estis sub lege 5:19 mani- 5:18 But if ye be led of the Spirit, ye are not under the law.

5:19 Now the works of the flesh are manifest, which are these; Adultery, fornication, uncleanness, lasciviousness,

festa autem sunt opera carnis quae sunt fornicatio inmunditia luxuria 5:20 idolorum seruitus ueneficia inimicitiae conten- 5:20 Idolatry, witchcraft, hatred, variance, emulations,

wrath, strife, seditions, heresies,

tiones aemulationes irae rixae dissensiones sectae 5:21 in- 5:21 Envyings, murders, drunkenness, revellings, and

such like: of the which I tell you before, as I have also told you in time past, that they which do such things shall

uidiae homicidia ebrietates comesationes et his similia quae not inherit the kingdom of God. praedico uobis sicut praedixi quoniam qui talia agunt reg-

num Dei non consequentur 5:22 fructus autem Spiritus est 5:22 But the fruit of the Spirit is love, joy, peace, longsuf-

fering, gentleness, goodness, faith,

caritas gaudium pax longanimitas bonitas benignitas 5:23 5:23 Meekness, temperance: against such there is no law. fides modestia continentia aduersus huiusmodi non est lex 5:24 qui autem sunt Christi carnem crucifixerunt cum uitiis 5:24 And they that are Christ’s have crucified the flesh

with the affections and lusts.

et concupiscentiis 5:25 si uiuimus spiritu spiritu et ambule- 5:25 If we live in the Spirit, let us also walk in the Spirit. mus 5:26 non efficiamur inanis gloriae cupidi inuicem pro- 5:26 Let us not be desirous of vain glory, provoking one

another, envying one another.

uocantes inuicem inuidentes

6:1 fratres et si praeoccupatus fuerit homo in aliquo delicto 6:1 Brethren, if a man be overtaken in a fault, ye which

are spiritual, restore such an one in the spirit of meekness; considering thyself, lest thou also be tempted.

uos qui spiritales estis huiusmodi instruite in spiritu lenitatis considerans te ipsum ne et tu tempteris 6:2 alter alterius 6:2 Bear ye one another’s burdens, and so fulfil the law of

Christ.

onera portate et sic adimplebitis legem Christi 6:3 nam si 6:3 For if a man think himself to be something, when he is

nothing, he deceiveth himself.

quis existimat se aliquid esse cum sit nihil ipse se seducit 6:4 opus autem suum probet unusquisque et sic in semet ipso 6:4 But let every man prove his own work, and then shall

he have rejoicing in himself alone, and not in another.

tantum gloriam habebit et non in altero 6:5 unusquisque 6:5 For every man shall bear his own burden. enim onus suum portabit 6:6 communicet autem is qui ca- 6:6 Let him that is taught in the word communicate unto

him that teacheth in all good things.

tecizatur uerbum ei qui se catecizat in omnibus bonis 6:7 6:7 Be not deceived; God is not mocked: for whatsoever a

man soweth, that shall he also reap.

nolite errare Deus non inridetur 6:8 quae enim seminaue- 6:8 For he that soweth to his flesh shall of the flesh reap

corruption; but he that soweth to the Spirit shall of the Spirit reap life everlasting.

rit homo haec et metet quoniam qui seminat in carne sua de carne et metet corruptionem qui autem seminat in spiritu

Biblia Sacra Vulgata

6:9 And let us not be weary in well doing: for in due sea- de spiritu metet uitam aeternam 6:9 bonum autem facientes

son we shall reap, if we faint not.

non deficiamus tempore enim suo metemus non deficientes

6:10 As we have therefore opportunity, let us do good unto 6:10 ergo dum tempus habemus operemur bonum ad omnes all men, especially unto them who are of the household of faith.

maxime autem ad domesticos fidei 6:11 uidete qualibus lit-

6:11 Ye see how large a letter I have written unto you with mine own hand.

6:12 As many as desire to make a fair shew in the flesh, teris scripsi uobis mea manu 6:12 quicumque uolunt pla-

they constrain you to be circumcised; only lest they should suffer persecution for the cross of Christ.

cere in carne hii cogunt uos circumcidi tantum ut crucis Ch-

6:13 For neither they themselves who are circumcised risti persecutionem non patiantur 6:13 neque enim qui cir-

keep the law; but desire to have you circumcised, that they may glory in your flesh.

cumciduntur legem custodiunt sed uolunt uos circumcidi ut

6:14 But God forbid that I should glory, save in the cross in carne uestra glorientur 6:14 mihi autem absit gloriari nisi

of our Lord Jesus Christ, by whom the world is crucified unto me, and I unto the world.

in cruce Domini nostri Iesu Christi per quem mihi mundus

6:15 For in Christ Jesus neither circumcision availeth any crucifixus est et ego mundo 6:15 in Christo enim Iesu neque thing, nor uncircumcision, but a new creature.

circumcisio aliquid ualet neque praeputium sed noua crea-

6:16 And as many as walk according to this rule, peace be tura 6:16 et quicumque hanc regulam secuti fuerint pax su-

on them, and mercy, and upon the Israel of God.

6:17 From henceforth let no man trouble me: for I bear in per illos et misericordia et super Israhel Dei 6:17 de cetero

my body the marks of the Lord Jesus.

nemo mihi molestus sit ego enim stigmata Iesu in corpore

6:18 Brethren, the grace of our Lord Jesus Christ be with meo porto 6:18 gratia Domini nostri Iesu Christi cum spi-

your spirit. Amen.

ritu uestro fratres amen

Epistola Pauli Ad Ephesios

1:1 Paulus apostolus Christi Iesu per uoluntatem Dei sanc- 1:1 Paul, an apostle of Jesus Christ by the will of God,

to the saints which are at Ephesus, and to the faithful in Christ Jesus:

tis omnibus qui sunt Ephesi et fidelibus in Christo Iesu 1:2

1:2 Grace be to you, and peace, from God our Father, and from the Lord Jesus Christ.

gratia uobis et pax a Deo Patre nostro et Domino Iesu Christo 1:3 benedictus Deus et Pater Domini nostri Iesu Ch- 1:3 Blessed be the God and Father of our Lord Jesus

Christ, who hath blessed us with all spiritual blessings in heavenly places in Christ:

risti qui benedixit nos in omni benedictione spiritali in caelestibus in Christo 1:4 sicut elegit nos in ipso ante mundi 1:4 According as he hath chosen us in him before the foun-

dation of the world, that we should be holy and without blame before him in love:

constitutionem ut essemus sancti et inmaculati in conspectu eius in caritate 1:5 qui praedestinauit nos in adoptionem 1:5 Having predestinated us unto the adoption of children

by Jesus Christ to himself, according to the good pleasure of his will,

filiorum per Iesum Christum in ipsum secundum propositum uoluntatis suae 1:6 in laudem gloriae gratiae suae in 1:6 To the praise of the glory of his grace, wherein he hath

made us accepted in the beloved.

qua gratificauit nos in dilecto 1:7 in quo habemus redeep- 1:7 In whom we have redemption through his blood, the

forgiveness of sins, according to the riches of his grace;

tionem per sanguinem eius remissionem peccatorum secundum diuitias gratiae eius 1:8 quae superabundauit in no- 1:8 Wherein he hath abounded toward us in all wisdom

and prudence;

bis in omni sapientia et prudentia 1:9 ut notum faceret no- 1:9 Having made known unto us the mystery of his will,

according to his good pleasure which he hath purposed in himself:

bis sacramentum uoluntatis suae secundum bonum placitum eius quod proposuit in eo 1:10 in dispensationem plenitu- 1:10 That in the dispensation of the fulness of times he

might gather together in one all things in Christ, both which are in heaven, and which are on earth; even in him:

dinis temporum instaurare omnia in Christo quae in caelis

et quae in terra sunt in ipso 1:11 in quo etiam sorte uo- 1:11 In whom also we have obtained an inheritance, being predestinated according to the purpose of him who worketh all things after the counsel of his own will:

cati sumus praedestinati secundum propositum eius qui omnia operatur secundum consilium uoluntatis suae 1:12 ut si- 1:12 That we should be to the praise of his glory, who first

trusted in Christ.

mus in laudem gloriae eius qui ante sperauimus in Christo 1:13 in quo et uos cum audissetis uerbum ueritatis euange- 1:13 In whom ye also trusted, after that ye heard the word

of truth, the gospel of your salvation: in whom also after that ye believed, ye were sealed with that holy Spirit of

lium salutis uestrae in quo et credentes signati estis Spiritu promise, promissionis Sancto 1:14 qui est pignus hereditatis nostrae 1:14 Which is the earnest of our inheritance until the re-

demption of the purchased possession, unto the praise of his glory.

in redeeptionem adquisitionis in laudem gloriae ipsius 1:15

1:15 Wherefore I also, after I heard of your faith in the Lord Jesus, and love unto all the saints,

propterea et ego audiens fidem uestram quae est in Domino Iesu et dilectionem in omnes sanctos 1:16 non cesso gra- 1:16 Cease not to give thanks for you, making mention of

you in my prayers;

tias agens pro uobis memoriam uestri faciens in orationibus meis 1:17 ut Deus Domini nostri Iesu Christi Pater glo- 1:17 That the God of our Lord Jesus Christ, the Father of

glory, may give unto you the spirit of wisdom and revelation in the knowledge of him:

riae det uobis spiritum sapientiae et reuelationis in agnitione eius 1:18 inluminatos oculos cordis uestri ut sciatis quae 1:18 The eyes of your understanding being enlightened;

that ye may know what is the hope of his calling, and what the riches of the glory of his inheritance in the saints,

sit spes uocationis eius quae diuitiae gloriae hereditatis eius in sanctis 1:19 et quae sit supereminens magnitudo uirtutis 1:19 And what is the exceeding greatness of his power

to us-ward who believe, according to the working of his mighty power,

Biblia Sacra Vulgata

eius in nos qui credidimus secundum operationem potentiae

1:20 Which he wrought in Christ, when he raised him from uirtutis eius 1:20 quam operatus est in Christo suscitans il-

the dead, and set him at his own right hand in the heavenly places,

lum a mortuis et constituens ad dexteram suam in caelestibus

1:21 Far above all principality, and power, and might, and 1:21 supra omnem principatum et potestatem et uirtutem et dominion, and every name that is named, not only in this world, but also in that which is to come:

dominationem et omne nomen quod nominatur non solum in

1:22 And hath put all things under his feet, and gave him hoc saeculo sed et in futuro 1:22 et omnia subiecit sub pe-

to be the head over all things to the church,

1:23 Which is his body, the fulness of him that filleth all dibus eius et ipsum dedit caput supra omnia ecclesiae 1:23

in all.

quae est corpus ipsius plenitudo eius qui omnia in omnibus adimpletur

2:1 And you hath he quickened, who were dead in tres- 2:1 et uos cum essetis mortui delictis et peccatis ue-

passes and sins;

2:2 Wherein in time past ye walked according to the stris 2:2 in quibus aliquando ambulastis secundum saecu-

course of this world, according to the prince of the power of the air, the spirit that now worketh in the children of

disobedience: lum mundi huius secundum principem potestatis aeris huius 2:3 Among whom also we all had our conversation in spiritus qui nunc operatur in filios diffidentiae 2:3 in quibus

times past in the lusts of our flesh, fulfilling the desires of the flesh and of the mind; and were by nature the children

of wrath, even as others. et nos omnes aliquando conuersati sumus in desideriis carnis

nostrae facientes uoluntates carnis et cogitationum et eramus

2:4 But God, who is rich in mercy, for his great love natura filii irae sicut et ceteri 2:4 Deus autem qui diues est

wherewith he loved us,

in misericordia propter nimiam caritatem suam qua dilexit

2:5 Even when we were dead in sins, hath quickened us nos 2:5 et cum essemus mortui peccatis conuiuificauit nos

together with Christ, (by grace ye are saved;)

2:6 And hath raised us up together, and made us sit to- Christo gratia estis saluati 2:6 et conresuscitauit et conse-

gether in heavenly places in Christ Jesus:

2:7 That in the ages to come he might shew the exceeding dere fecit in caelestibus in Christo Iesu 2:7 ut ostenderet in

riches of his grace in his kindness toward us through Christ Jesus.

saeculis superuenientibus abundantes diuitias gratiae suae in

2:8 For by grace are ye saved through faith; and that not bonitate super nos in Christo Iesu 2:8 gratia enim estis sal-

of yourselves: it is the gift of God:

2:9 Not of works, lest any man should boast. uati per fidem et hoc non ex uobis Dei enim donum est 2:9 2:10 For we are his workmanship, created in Christ Jesus non ex operibus ut ne quis glorietur 2:10 ipsius enim sumus

unto good works, which God hath before ordained that we should walk in them.

factura creati in Christo Iesu in operibus bonis quae praepa-

2:11 Wherefore remember, that ye being in time past Gen- rauit Deus ut in illis ambulemus 2:11 propter quod memotiles in the flesh, who are called Uncircumcision by that which is called the Circumcision in the flesh made by

hands; res estote quod aliquando uos gentes in carne qui dicimini

praeputium ab ea quae dicitur circumcisio in carne manu-

2:12 That at that time ye were without Christ, being aliens facta 2:12 quia eratis illo in tempore sine Christo alienati a

from the commonwealth of Israel, and strangers from the covenants of promise, having no hope, and without God in

the world: conuersatione Israhel et hospites testamentorum promissio-2:13 But now in Christ Jesus ye who sometimes were far nis spem non habentes et sine Deo in mundo 2:13 nunc au-

off are made nigh by the blood of Christ.

tem in Christo Iesu uos qui aliquando eratis longe facti estis

2:14 For he is our peace, who hath made both one, and prope in sanguine Christi 2:14 ipse est enim pax nostra qui

hath broken down the middle wall of partition between us;

Epistola Pauli Ad Ephesios

fecit utraque unum et medium parietem maceriae soluens inimicitiam in carne sua 2:15 legem mandatorum decretis 2:15 Having abolished in his flesh the enmity, even the law

of commandments contained in ordinances; for to make in himself of twain one new man, so making peace;

euacuans ut duos condat in semet ipsum in unum nouum hominem faciens pacem 2:16 et reconciliet ambos in uno cor- 2:16 And that he might reconcile both unto God in one

body by the cross, having slain the enmity thereby:

pore Deo per crucem interficiens inimicitiam in semet ipso 2:17 et ueniens euangelizauit pacem uobis qui longe fuistis et 2:17 And came and preached peace to you which were afar

off, and to them that were nigh.

pacem his qui prope 2:18 quoniam per ipsum habemus ac- 2:18 For through him we both have access by one Spirit

unto the Father.

cessum ambo in uno Spiritu ad Patrem 2:19 ergo iam non 2:19 Now therefore ye are no more strangers and foreigners, but fellowcitizens with the saints, and of the household of God;

estis hospites et aduenae sed estis ciues sanctorum et domestici Dei 2:20 superaedificati super fundamentum apostolo- 2:20 And are built upon the foundation of the apostles

and prophets, Jesus Christ himself being the chief corner stone;

rum et prophetarum ipso summo angulari lapide Christo Iesu 2:21 in quo omnis aedificatio constructa crescit in templum 2:21 In whom all the building fitly framed together

groweth unto an holy temple in the Lord:

sanctum in Domino 2:22 in quo et uos coaedificamini in 2:22 In whom ye also are builded together for an habita-

tion of God through the Spirit.

habitaculum Dei in Spiritu

3:1 huius rei gratia ego Paulus uinctus Christi Iesu pro uo- 3:1 For this cause I Paul, the prisoner of Jesus Christ for

you Gentiles,

bis gentibus 3:2 si tamen audistis dispensationem gratiae 3:2 If ye have heard of the dispensation of the grace of

God which is given me to you-ward:

Dei quae data est mihi in uobis 3:3 quoniam secundum re- 3:3 How that by revelation he made known unto me the

mystery; (as I wrote afore in few words,

uelationem notum mihi factum est sacramentum sicut supra scripsi in breui 3:4 prout potestis legentes intellegere pru- 3:4 Whereby, when ye read, ye may understand my knowl-

edge in the mystery of Christ)

dentiam meam in mysterio Christi 3:5 quod aliis genera- 3:5 Which in other ages was not made known unto the

sons of men, as it is now revealed unto his holy apostles and prophets by the Spirit;

tionibus non est agnitum filiis hominum sicuti nunc reuelatum est sanctis apostolis eius et prophetis in Spiritu 3:6 esse 3:6 That the Gentiles should be fellowheirs, and of the

same body, and partakers of his promise in Christ by the gospel:

gentes coheredes et concorporales et conparticipes promissionis in Christo Iesu per euangelium 3:7 cuius factus sum 3:7 Whereof I was made a minister, according to the gift of

the grace of God given unto me by the effectual working of his power.

minister secundum donum gratiae Dei quae data est mihi secundum operationem uirtutis eius 3:8 mihi omnium sancto- 3:8 Unto me, who am less than the least of all saints, is

this grace given, that I should preach among the Gentiles the unsearchable riches of Christ;

rum minimo data est gratia haec in gentibus euangelizare ininuestigabiles diuitias Christi 3:9 et inluminare omnes quae 3:9 And to make all men see what is the fellowship of the

mystery, which from the beginning of the world hath been hid in God, who created all things by Jesus Christ:

sit dispensatio sacramenti absconditi a saeculis in Deo qui omnia creauit 3:10 ut innotescat principibus et potestatibus 3:10 To the intent that now unto the principalities and pow-

ers in heavenly places might be known by the church the manifold wisdom of God,

in caelestibus per ecclesiam multiformis sapientia Dei 3:11

3:11 According to the eternal purpose which he purposed in Christ Jesus our Lord:

secundum praefinitionem saeculorum quam fecit in Christo Iesu Domino nostro 3:12 in quo habemus fiduciam et ac- 3:12 In whom we have boldness and access with confi-

dence by the faith of him.

3:13 Wherefore I desire that ye faint not at my tribulations cessum in confidentia per fidem eius 3:13 propter quod peto for you, which is your glory.

ne deficiatis in tribulationibus meis pro uobis quae est glo-

3:14 For this cause I bow my knees unto the Father of our ria uestra 3:14 huius rei gratia flecto genua mea ad Patrem

Lord Jesus Christ,

3:15 Of whom the whole family in heaven and earth is Domini nostri Iesu Christi 3:15 ex quo omnis paternitas in

named,

3:16 That he would grant you, according to the riches of caelis et in terra nominatur 3:16 ut det uobis secundum di-

his glory, to be strengthened with might by his Spirit in the inner man;

uitias gloriae suae uirtute corroborari per Spiritum eius in

3:17 That Christ may dwell in your hearts by faith; that ye, interiore homine 3:17 habitare Christum per fidem in corbeing rooted and grounded in love,

3:18 May be able to comprehend with all saints what is dibus uestris in caritate radicati et fundati 3:18 ut possitis

the breadth, and length, and depth, and height;

conprehendere cum omnibus sanctis quae sit latitudo et lon-

3:19 And to know the love of Christ, which passeth knowl- gitudo et sublimitas et profundum 3:19 scire etiam supere-

edge, that ye might be filled with all the fulness of God.

minentem scientiae caritatem Christi ut impleamini in om-

3:20 Now unto him that is able to do exceeding abundantly nem plenitudinem Dei 3:20 ei autem qui potens est omnia above all that we ask or think, according to the power that worketh in us,

facere superabundanter quam petimus aut intellegimus se-

3:21 Unto him be glory in the church by Christ Jesus cundum uirtutem quae operatur in nobis 3:21 ipsi gloria in

throughout all ages, world without end. Amen.

ecclesia et in Christo Iesu in omnes generationes saeculi saeculorum amen

4:1 I therefore, the prisoner of the Lord, beseech you that 4:1 obsecro itaque uos ego uinctus in Domino ut digne amye walk worthy of the vocation wherewith ye are called,

4:2 With all lowliness and meekness, with longsuffering, buletis uocatione qua uocati estis 4:2 cum omni humilitate

forbearing one another in love;

et mansuetudine cum patientia subportantes inuicem in ca-

4:3 Endeavouring to keep the unity of the Spirit in the ritate 4:3 solliciti seruare unitatem spiritus in uinculo pacis

bond of peace.

4:4 There is one body, and one Spirit, even as ye are called 4:4 unum corpus et unus spiritus sicut uocati estis in una spe

in one hope of your calling;

4:5 One Lord, one faith, one baptism, uocationis uestrae 4:5 unus Dominus una fides unum ba-4:6 One God and Father of all, who is above all, and ptisma 4:6 unus Deus et Pater omnium qui super omnes

through all, and in you all.

4:7 But unto every one of us is given grace according to et per omnia et in omnibus nobis 4:7 unicuique autem no-

the measure of the gift of Christ.

strum data est gratia secundum mensuram donationis Christi

4:8 Wherefore he saith, When he ascended up on high, he 4:8 propter quod dicit ascendens in altum captiuam duxit ca-

led captivity captive, and gave gifts unto men.

4:9 (Now that he ascended, what is it but that he also de- ptiuitatem dedit dona hominibus 4:9 quod autem ascendit

scended first into the lower parts of the earth?

quid est nisi quia et descendit primum in inferiores partes

4:10 He that descended is the same also that ascended up terrae 4:10 qui descendit ipse est et qui ascendit super om-

far above all heavens, that he might fill all things.) 4:11 And he gave some, apostles; and some, prophets; and nes caelos ut impleret omnia 4:11 et ipse dedit quosdam some, evangelists; and some, pastors and teachers;

quidem apostolos quosdam autem prophetas alios uero euan-

4:12 For the perfecting of the saints, for the work of the gelistas alios autem pastores et doctores 4:12 ad consum-

ministry, for the edifying of the body of Christ:

mationem sanctorum in opus ministerii in aedificationem

Epistola Pauli Ad Ephesios

corporis Christi 4:13 donec occurramus omnes in unitatem 4:13 Till we all come in the unity of the faith, and of the

knowledge of the Son of God, unto a perfect man, unto the measure of the stature of the fulness of Christ:

fidei et agnitionis Filii Dei in uirum perfectum in mensuram aetatis plenitudinis Christi 4:14 ut iam non simus paruuli 4:14 That we henceforth be no more children, tossed to

and fro, and carried about with every wind of doctrine, by the sleight of men, and cunning craftiness, whereby they

fluctuantes et circumferamur omni uento doctrinae in nequi- lie in wait to deceive; tia hominum in astutia ad circumuentionem erroris 4:15 ue- 4:15 But speaking the truth in love, may grow up into him

in all things, which is the head, even Christ:

ritatem autem facientes in caritate crescamus in illo per omnia qui est caput Christus 4:16 ex quo totum corpus con- 4:16 From whom the whole body fitly joined together and

compacted by that which every joint supplieth, according to the effectual working in the measure of every part, maketh increase of the body unto the edifying of itself in love.

pactum et conexum per omnem iuncturam subministrationis secundum operationem in mensuram uniuscuiusque membri augmentum corporis facit in aedificationem sui in caritate 4:17 hoc igitur dico et testificor in Domino ut iam non 4:17 This I say therefore, and testify in the Lord, that ye

henceforth walk not as other Gentiles walk, in the vanity of their mind,

ambuletis sicut gentes ambulant in uanitate sensus sui 4:18

4:18 Having the understanding darkened, being alienated from the life of God through the ignorance that is in them, because of the blindness of their heart:

tenebris obscuratum habentes intellectum alienati a uita Dei per ignorantiam quae est in illis propter caecitatem cordis ipsorum 4:19 qui desperantes semet ipsos tradiderunt inpu- 4:19 Who being past feeling have given themselves over

unto lasciviousness, to work all uncleanness with greediness.

dicitiae in operationem inmunditiae omnis in auaritia 4:20

4:20 But ye have not so learned Christ;

uos autem non ita didicistis Christum 4:21 si tamen illum 4:21 If so be that ye have heard him, and have been taught

by him, as the truth is in Jesus:

audistis et in ipso edocti estis sicut est ueritas in Iesu 4:22 4:22 That ye put off concerning the former conversation

the old man, which is corrupt according to the deceitful lusts;

deponere uos secundum pristinam conuersationem ueterem hominem qui corrumpitur secundum desideria erroris 4:23 4:23 And be renewed in the spirit of your mind; renouamini autem spiritu mentis uestrae 4:24 et induite no- 4:24 And that ye put on the new man, which after God is

created in righteousness and true holiness.

uum hominem qui secundum Deum creatus est in iustitia et sanctitate ueritatis 4:25 propter quod deponentes men- 4:25 Wherefore putting away lying, speak every man truth

with his neighbour: for we are members one of another.

dacium loquimini ueritatem unusquisque cum proximo suo quoniam sumus inuicem membra 4:26 irascimini et nolite 4:26 Be ye angry, and sin not: let not the sun go down

upon your wrath:

peccare sol non occidat super iracundiam uestram 4:27 no- 4:27 Neither give place to the devil. lite locum dare diabolo 4:28 qui furabatur iam non furetur 4:28 Let him that stole steal no more: but rather let him

labour, working with his hands the thing which is good, that he may have to give to him that needeth.

magis autem laboret operando manibus quod bonum est ut habeat unde tribuat necessitatem patienti 4:29 omnis sermo 4:29 Let no corrupt communication proceed out of your

mouth, but that which is good to the use of edifying, that it may minister grace unto the hearers.

malus ex ore uestro non procedat sed si quis bonus ad aedificationem oportunitatis ut det gratiam audientibus 4:30 et 4:30 And grieve not the holy Spirit of God, whereby ye

are sealed unto the day of redemption.

nolite contristare Spiritum Sanctum Dei in quo signati estis in die redeeptionis 4:31 omnis amaritudo et ira et indigna- 4:31 Let all bitterness, and wrath, and anger, and clamour,

and evil speaking, be put away from you, with all malice:

Biblia Sacra Vulgata

tio et clamor et blasphemia tollatur a uobis cum omni malitia

4:32 And be ye kind one to another, tenderhearted, forgiv- 4:32 estote autem inuicem benigni misericordes donantes in-

ing one another, even as God for Christ’s sake hath forgiven you.

uicem sicut et Deus in Christo donauit nobis

5:1 Be ye therefore followers of God, as dear children; 5:1 estote ergo imitatores Dei sicut filii carissimi 5:2 et

5:2 And walk in love, as Christ also hath loved us, and hath given himself for us an offering and a sacrifice to God for a sweetsmelling savour.

ambulate in dilectione sicut et Christus dilexit nos et tradidit se ipsum pro nobis oblationem et hostiam Deo in odo-

5:3 But fornication, and all uncleanness, or covetousness, rem suauitatis 5:3 fornicatio autem et omnis inmunditia aut let it not be once named among you, as becometh saints;

5:4 Neither filthiness, nor foolish talking, nor jesting, auaritia nec nominetur in uobis sicut decet sanctos 5:4 aut

which are not convenient: but rather giving of thanks.

turpitudo aut stultiloquium aut scurrilitas quae ad rem non

5:5 For this ye know, that no whoremonger, nor unclean pertinent sed magis gratiarum actio 5:5 hoc enim scitote in-

person, nor covetous man, who is an idolater, hath any inheritance in the kingdom of Christ and of God.

tellegentes quod omnis fornicator aut inmundus aut auarus quod est idolorum seruitus non habet hereditatem in regno

5:6 Let no man deceive you with vain words: for because Christi et Dei 5:6 nemo uos seducat inanibus uerbis propter

of these things cometh the wrath of God upon the children of disobedience.

haec enim uenit ira Dei in filios diffidentiae 5:7 nolite ergo

5:7 Be not ye therefore partakers with them.

5:8 For ye were sometimes darkness, but now are ye light effici participes eorum 5:8 eratis enim aliquando tenebrae

in the Lord: walk as children of light: 5:9 (For the fruit of the Spirit is in all goodness and righ- nunc autem lux in Domino ut filii lucis ambulate 5:9 fructeousness and truth;)

5:10 Proving what is acceptable unto the Lord. tus enim lucis est in omni bonitate et iustitia et ueritate 5:10 5:11 And have no fellowship with the unfruitful works of probantes quid sit beneplacitum Deo 5:11 et nolite commu-

darkness, but rather reprove them.

nicare operibus infructuosis tenebrarum magis autem et re-

5:12 For it is a shame even to speak of those things which darguite 5:12 quae enim in occulto fiunt ab ipsis turpe est

are done of them in secret. 5:13 But all things that are reproved are made manifest by et dicere 5:13 omnia autem quae arguuntur a lumine manithe light: for whatsoever doth make manifest is light.

5:14 Wherefore he saith, Awake thou that sleepest, and festantur omne enim quod manifestatur lumen est 5:14 pro-

arise from the dead, and Christ shall give thee light.

pter quod dicit surge qui dormis et exsurge a mortuis et in-

5:15 See then that ye walk circumspectly, not as fools, but luminabit tibi Christus 5:15 uidete itaque fratres quomodo

as wise,

5:16 Redeeming the time, because the days are evil. caute ambuletis non quasi insipientes sed ut sapientes 5:16 5:17 Wherefore be ye not unwise, but understanding what rediientes tempus quoniam dies mali sunt 5:17 propterea

the will of the Lord is.

nolite fieri inprudentes sed intellegentes quae sit uoluntas

5:18 And be not drunk with wine, wherein is excess; but Domini 5:18 et nolite inebriari uino in quo est luxuria sed

be filled with the Spirit;

5:19 Speaking to yourselves in psalms and hymns and implemini Spiritu 5:19 loquentes uobismet ipsis in psalmis

spiritual songs, singing and making melody in your heart to the Lord;

et hymnis et canticis spiritalibus cantantes et psallentes in

5:20 Giving thanks always for all things unto God and the cordibus uestris Domino 5:20 gratias agentes semper pro

Father in the name of our Lord Jesus Christ;

omnibus in nomine Domini nostri Iesu Christi Deo et Patri

5:21 Submitting yourselves one to another in the fear of 5:21 subiecti inuicem in timore Christi 5:22 mulieres uiris

God. 5:22 Wives, submit yourselves unto your own husbands, as unto the Lord.

Epistola Pauli Ad Ephesios

suis subditae sint sicut Domino 5:23 quoniam uir caput est 5:23 For the husband is the head of the wife, even as Christ

is the head of the church: and he is the saviour of the body.

mulieris sicut Christus caput est ecclesiae ipse saluator corporis 5:24 sed ut ecclesia subiecta est Christo ita et mulieres 5:24 Therefore as the church is subject unto Christ, so let

the wives be to their own husbands in every thing.

uiris suis in omnibus 5:25 uiri diligite uxores sicut et Chri- 5:25 Husbands, love your wives, even as Christ also loved

the church, and gave himself for it;

stus dilexit ecclesiam et se ipsum tradidit pro ea 5:26 ut 5:26 That he might sanctify and cleanse it with the wash-

ing of water by the word,

illam sanctificaret mundans lauacro aquae in uerbo 5:27 ut 5:27 That he might present it to himself a glorious church,

not having spot, or wrinkle, or any such thing; but that it should be holy and without blemish.

exhiberet ipse sibi gloriosam ecclesiam non habentem maculam aut rugam aut aliquid eiusmodi sed ut sit sancta et inmaculata 5:28 ita et uiri debent diligere uxores suas ut 5:28 So ought men to love their wives as their own bodies.

He that loveth his wife loveth himself.

corpora sua qui suam uxorem diligit se ipsum diligit 5:29 5:29 For no man ever yet hated his own flesh; but nour-

isheth and cherisheth it, even as the Lord the church:

nemo enim umquam carnem suam odio habuit sed nutrit et fouet eam sicut et Christus ecclesiam 5:30 quia membra su- 5:30 For we are members of his body, of his flesh, and of

his bones.

mus corporis eius de carne eius et de ossibus eius 5:31 pro- 5:31 For this cause shall a man leave his father and mother,

and shall be joined unto his wife, and they two shall be one flesh.

pter hoc relinquet homo patrem et matrem suam et adherebit

uxori suae et erunt duo in carne una 5:32 sacramentum hoc 5:32 This is a great mystery: but I speak concerning Christ

and the church.

magnum est ego autem dico in Christo et in ecclesia 5:33 5:33 Nevertheless let every one of you in particular so love his wife even as himself; and the wife see that she reverence her husband.

uerumtamen et uos singuli unusquisque suam uxorem sicut se ipsum diligat uxor autem ut timeat uirum

6:1 filii oboedite parentibus uestris in Domino hoc enim 6:1 Children, obey your parents in the Lord: for this is

right.

est iustum 6:2 honora patrem tuum et matrem quod est 6:2 Honour thy father and mother; which is the first com-

mandment with promise;

mandatum primum in promissione 6:3 ut bene sit tibi et sis 6:3 That it may be well with thee, and thou mayest live

long on the earth.

longeuus super terram 6:4 et patres nolite ad iracundiam 6:4 And, ye fathers, provoke not your children to wrath:

but bring them up in the nurture and admonition of the Lord.

prouocare filios uestros sed educate illos in disciplina et correptione Domini 6:5 serui oboedite dominis carnalibus cum 6:5 Servants, be obedient to them that are your masters ac-

cording to the flesh, with fear and trembling, in singleness of your heart, as unto Christ;

timore et tremore in simplicitate cordis uestri sicut Christo 6:6 non ad oculum seruientes quasi hominibus placentes sed 6:6 Not with eyeservice, as menpleasers; but as the ser-

vants of Christ, doing the will of God from the heart;

ut serui Christi facientes uoluntatem Dei ex animo 6:7 cum 6:7 With good will doing service, as to the Lord, and not

to men:

bona uoluntate seruientes sicut Domino et non hominibus 6:8 scientes quoniam unusquisque quodcumque fecerit bo- 6:8 Knowing that whatsoever good thing any man doeth,

the same shall he receive of the Lord, whether he be bond or free.

num hoc percipiet a Domino siue seruus siue liber 6:9 et

6:9 And, ye masters, do the same things unto them, forbearing threatening: knowing that your Master also is in heaven; neither is there respect of persons with him.

domini eadem facite illis remittentes minas scientes quia et illorum et uester Dominus est in caelis et personarum acce-

ptio non est apud eum 6:10 de cetero fratres confortamini 6:10 Finally, my brethren, be strong in the Lord, and in the power of his might.

Biblia Sacra Vulgata

6:11 Put on the whole armour of God, that ye may be able in Domino et in potentia uirtutis eius 6:11 induite uos arma

to stand against the wiles of the devil.

6:12 For we wrestle not against flesh and blood, but Dei ut possitis stare aduersus insidias diaboli 6:12 quia non

against principalities, against powers, against the rulers of the darkness of this world, against spiritual wickedness in

high places. est nobis conluctatio aduersus carnem et sanguinem sed ad-

uersus principes et potestates aduersus mundi rectores tene-

6:13 Wherefore take unto you the whole armour of God, brarum harum contra spiritalia nequitiae in caelestibus 6:13

that ye may be able to withstand in the evil day, and having done all, to stand.

propterea accipite armaturam Dei ut possitis resistere in die

6:14 Stand therefore, having your loins girt about with malo et omnibus perfectis stare 6:14 state ergo succincti

truth, and having on the breastplate of righteousness;

6:15 And your feet shod with the preparation of the gospel lumbos uestros in ueritate et induti loricam iustitiae 6:15 et

of peace;

6:16 Above all, taking the shield of faith, wherewith ye calciati pedes in praeparatione euangelii pacis 6:16 in om-

shall be able to quench all the fiery darts of the wicked.

nibus sumentes scutum fidei in quo possitis omnia tela ne-

6:17 And take the helmet of salvation, and the sword of quissimi ignea extinguere 6:17 et galeam salutis adsumite

the Spirit, which is the word of God: 6:18 Praying always with all prayer and supplication in et gladium Spiritus quod est uerbum Dei 6:18 per omnem the Spirit, and watching thereunto with all perseverance and supplication for all saints;

orationem et obsecrationem orantes omni tempore in Spiritu et in ipso uigilantes in omni instantia et obsecratione pro

6:19 And for me, that utterance may be given unto me, that omnibus sanctis 6:19 et pro me ut detur mihi sermo in aper-

I may open my mouth boldly, to make known the mystery of the gospel,

tione oris mei cum fiducia notum facere mysterium euangelii

6:20 For which I am an ambassador in bonds: that therein 6:20 pro quo legatione fungor in catena ita ut in ipso audeam

I may speak boldly, as I ought to speak.

6:21 But that ye also may know my affairs, and how I do, prout oportet me loqui 6:21 ut autem et uos sciatis quae

Tychicus, a beloved brother and faithful minister in the Lord, shall make known to you all things:

circa me sunt quid agam omnia nota uobis faciet Tychicus

6:22 Whom I have sent unto you for the same purpose, carissimus frater et fidelis minister in Domino 6:22 quem

that ye might know our affairs, and that he might comfort your hearts.

misi ad uos in hoc ipsum ut cognoscatis quae circa nos sunt

6:23 Peace be to the brethren, and love with faith, from et consoletur corda uestra 6:23 pax fratribus et caritas cum

God the Father and the Lord Jesus Christ. 6:24 Grace be with all them that love our Lord Jesus Christ fide a Deo Patre et Domino Iesu Christo 6:24 gratia cum in sincerity. Amen.

omnibus qui diligunt Dominum nostrum Iesum Christum in incorruptione

Epistola Pauli Ad Philippenses

1:1 Paulus et Timotheus serui Iesu Christi omnibus sanctis 1:1 Paul and Timotheus, the servants of Jesus Christ, to all

the saints in Christ Jesus which are at Philippi, with the bishops and deacons:

in Christo Iesu qui sunt Philippis cum episcopis et diaconis 1:2 gratia uobis et pax a Deo Patre nostro et Domino Iesu 1:2 Grace be unto you, and peace, from God our Father,

and from the Lord Jesus Christ.

Christo 1:3 gratias ago Deo meo in omni memoria uestri 1:3 I thank my God upon every remembrance of you, 1:4 semper in cunctis orationibus meis pro omnibus uobis 1:4 Always in every prayer of mine for you all making

request with joy,

cum gaudio deprecationem faciens 1:5 super communica- 1:5 For your fellowship in the gospel from the first day

until now;

tione uestra in euangelio a prima die usque nunc 1:6 con- 1:6 Being confident of this very thing, that he which hath

begun a good work in you will perform it until the day of Jesus Christ:

fidens hoc ipsum quia qui coepit in uobis opus bonum perficiet usque in diem Christi Iesu 1:7 sicut est mihi iustum 1:7 Even as it is meet for me to think this of you all, be-

cause I have you in my heart; inasmuch as both in my bonds, and in the defence and confirmation of the gospel,

hoc sentire pro omnibus uobis eo quod habeam in corde uos ye all are partakers of my grace. et in uinculis meis et in defensione et confirmatione euangelii socios gaudii mei omnes uos esse 1:8 testis enim mihi 1:8 For God is my record, how greatly I long after you all

in the bowels of Jesus Christ.

est Deus quomodo cupiam omnes uos in uisceribus Christi Iesu 1:9 et hoc oro ut caritas uestra magis ac magis abun- 1:9 And this I pray, that your love may abound yet more

and more in knowledge and in all judgment;

det in scientia et omni sensu 1:10 ut probetis potiora ut sitis 1:10 That ye may approve things that are excellent; that ye

may be sincere and without offence till the day of Christ.

sinceres et sine offensa in diem Christi 1:11 repleti fructu 1:11 Being filled with the fruits of righteousness, which

are by Jesus Christ, unto the glory and praise of God.

iustitiae per Christum Iesum in gloriam et laudem Dei 1:12 1:12 But I would ye should understand, brethren, that the

things which happened unto me have fallen out rather unto the furtherance of the gospel;

scire autem uos uolo fratres quia quae circa me sunt magis ad profectum uenerunt euangelii 1:13 ita ut uincula mea ma- 1:13 So that my bonds in Christ are manifest in all the

palace, and in all other places;

nifesta fierent in Christo in omni praetorio et in ceteris omnibus 1:14 et plures e fratribus in Domino confidentes uin- 1:14 And many of the brethren in the Lord, waxing confi-

dent by my bonds, are much more bold to speak the word without fear.

culis meis abundantius audere sine timore uerbum Dei loqui 1:15 quidam quidem et propter inuidiam et contentionem qui- 1:15 Some indeed preach Christ even of envy and strife;

and some also of good will:

dam autem et propter bonam uoluntatem Christum praedicant 1:16 quidam ex caritate scientes quoniam in defensio- 1:16 The one preach Christ of contention, not sincerely,

supposing to add affliction to my bonds:

nem euangelii positus sum 1:17 quidam autem ex conten- 1:17 But the other of love, knowing that I am set for the

defence of the gospel.

tione Christum adnuntiant non sincere existimantes pressu-

ram se suscitare uinculis meis 1:18 quid enim dum omni 1:18 What then? notwithstanding, every way, whether in pretence, or in truth, Christ is preached; and I therein do rejoice, yea, and will rejoice.

modo siue per occasionem siue per ueritatem Christus ad-

nuntiatur et in hoc gaudeo sed et gaudebo 1:19 scio enim 1:19 For I know that this shall turn to my salvation through

your prayer, and the supply of the Spirit of Jesus Christ,

quia hoc mihi proueniet in salutem per uestram orationem et subministrationem Spiritus Iesu Christi 1:20 secundum 1:20 According to my earnest expectation and my hope,

that in nothing I shall be ashamed, but that with all boldness, as always, so now also Christ shall be magnified in

expectationem et spem meam quia in nullo confundar sed in my body, whether it be by life, or by death.

Biblia Sacra Vulgata

omni fiducia sicut semper et nunc magnificabitur Christus in

1:21 For to me to live is Christ, and to die is gain. corpore meo siue per uitam siue per mortem 1:21 mihi enim

1:22 But if I live in the flesh, this is the fruit of my labour: uiuere Christus est et mori lucrum 1:22 quod si uiuere in yet what I shall choose I wot not.

1:23 For I am in a strait betwixt two, having a desire to carne hic mihi fructus operis est et quid eligam ignoro 1:23

depart, and to be with Christ; which is far better:

coartor autem e duobus desiderium habens dissolui et cum

1:24 Nevertheless to abide in the flesh is more needful for Christo esse multo magis melius 1:24 permanere autem in you.

1:25 And having this confidence, I know that I shall abide carne magis necessarium est propter uos 1:25 et hoc confi-

and continue with you all for your furtherance and joy of faith;

dens scio quia manebo et permanebo omnibus uobis ad pro-

1:26 That your rejoicing may be more abundant in Jesus fectum uestrum et gaudium fidei 1:26 ut gratulatio uestra

Christ for me by my coming to you again.

abundet in Christo Iesu in me per meum aduentum iterum

1:27 Only let your conversation be as it becometh the ad uos 1:27 tantum digne euangelio Christi conuersamini ut

gospel of Christ: that whether I come and see you, or else be absent, I may hear of your affairs, that ye stand fast in one spirit, with one mind striving together for the faith of the gospel;

siue cum uenero et uidero uos siue absens audiam de uobis quia stetistis uno spiritu unianimes conlaborantes fide euan-

1:28 And in nothing terrified by your adversaries: which gelii 1:28 et in nullo terreamini ab aduersariis quae est illis

is to them an evident token of perdition, but to you of salvation, and that of God.

causa perditionis uobis autem salutis et hoc a Deo 1:29 quia

1:29 For unto you it is given in the behalf of Christ, not only to believe on him, but also to suffer for his sake;

uobis donatum est pro Christo non solum ut in eum creda-

1:30 Having the same conflict which ye saw in me, and tis sed ut etiam pro illo patiamini 1:30 eundem certamen

now hear to be in me.

habentes qualem et uidistis in me et nunc audistis de me

2:1 If there be therefore any consolation in Christ, if any 2:1 si qua ergo consolatio in Christo si quod solacium cacomfort of love, if any fellowship of the Spirit, if any bowels and mercies,

ritatis si qua societas spiritus si quid uiscera et miseratio-

2:2 Fulfil ye my joy, that ye be likeminded, having the nes 2:2 implete gaudium meum ut idem sapiatis eandem same love, being of one accord, of one mind.

2:3 Let nothing be done through strife or vainglory; but in caritatem habentes unianimes id ipsum sentientes 2:3 nihil

lowliness of mind let each esteem other better than themselves.

per contentionem neque per inanem gloriam sed in humi-

2:4 Look not every man on his own things, but every man litate superiores sibi inuicem arbitrantes 2:4 non quae sua

also on the things of others. 2:5 Let this mind be in you, which was also in Christ Jesus: sunt singuli considerantes sed et ea quae aliorum 2:5 hoc 2:6 Who, being in the form of God, thought it not robbery enim sentite in uobis quod et in Christo Iesu 2:6 qui cum to be equal with God:

in forma Dei esset non rapinam arbitratus est esse se aequa-

2:7 But made himself of no reputation, and took upon him lem Deo 2:7 sed semet ipsum exinaniuit formam serui acthe form of a servant, and was made in the likeness of men:

cipiens in similitudinem hominum factus et habitu inuentus

2:8 And being found in fashion as a man, he humbled him- ut homo 2:8 humiliauit semet ipsum factus oboediens us-

self, and became obedient unto death, even the death of the cross.

que ad mortem mortem autem crucis 2:9 propter quod et

2:9 Wherefore God also hath highly exalted him, and given him a name which is above every name:

Deus illum exaltauit et donauit illi nomen super omne no-

2:10 That at the name of Jesus every knee should bow, of men 2:10 ut in nomine Iesu omne genu flectat caelestium things in heaven, and things in earth, and things under the earth;

Epistola Pauli Ad Philippenses

et terrestrium et infernorum 2:11 et omnis lingua confite- 2:11 And that every tongue should confess that Jesus

Christ is Lord, to the glory of God the Father.

atur quia Dominus Iesus Christus in gloria est Dei Patris 2:12 itaque carissimi mei sicut semper oboedistis non ut in 2:12 Wherefore, my beloved, as ye have always obeyed,

not as in my presence only, but now much more in my absence, work out your own salvation with fear and trem-

praesentia mei tantum sed multo magis nunc in absentia mea bling. cum metu et tremore uestram salutem operamini 2:13 Deus 2:13 For it is God which worketh in you both to will and

to do of his good pleasure.

est enim qui operatur in uobis et uelle et perficere pro bona uoluntate 2:14 omnia autem facite sine murmurationibus et 2:14 Do all things without murmurings and disputings: haesitationibus 2:15 ut sitis sine querella et simplices filii 2:15 That ye may be blameless and harmless, the sons of

God, without rebuke, in the midst of a crooked and perverse nation, among whom ye shine as lights in the world;

Dei sine reprehensione in medio nationis prauae et peruersae inter quos lucetis sicut luminaria in mundo 2:16 uer- 2:16 Holding forth the word of life; that I may rejoice

in the day of Christ, that I have not run in vain, neither laboured in vain.

bum uitae continentes ad gloriam meam in die Christi quia

non in uacuum cucurri neque in uacuum laboraui 2:17 sed 2:17 Yea, and if I be offered upon the sacrifice and service

of your faith, I joy, and rejoice with you all.

et si immolor supra sacrificium et obsequium fidei uestrae gaudeo et congratulor omnibus uobis 2:18 id ipsum autem 2:18 For the same cause also do ye joy, and rejoice with

me.

et uos gaudete et congratulamini mihi 2:19 spero autem in 2:19 But I trust in the Lord Jesus to send Timotheus shortly unto you, that I also may be of good comfort, when I know your state.

Domino Iesu Timotheum cito me mittere ad uos ut et ego bono animo sim cognitis quae circa uos sunt 2:20 neminem 2:20 For I have no man likeminded, who will naturally

care for your state.

enim habeo tam unianimem qui sincera affectione pro uo-

bis sollicitus sit 2:21 omnes enim sua quaerunt non quae 2:21 For all seek their own, not the things which are Jesus

Christ’s.

sunt Christi Iesu 2:22 experimentum autem eius cognos- 2:22 But ye know the proof of him, that, as a son with the

father, he hath served with me in the gospel.

cite quoniam sicut patri filius mecum seruiuit in euangelium 2:23 hunc igitur spero me mittere mox ut uidero quae circa 2:23 Him therefore I hope to send presently, so soon as I

shall see how it will go with me.

me sunt 2:24 confido autem in Domino quoniam et ipse ue- 2:24 But I trust in the Lord that I also myself shall come

shortly.

niam ad uos cito 2:25 necessarium autem existimaui Epa- 2:25 Yet I supposed it necessary to send to you Epaphroditus, my brother, and companion in labour, and fellowsoldier, but your messenger, and he that ministered to my

froditum fratrem et cooperatorem et commilitonem meum wants. uestrum autem apostolum et ministrum necessitatis meae

mittere ad uos 2:26 quoniam quidem omnes uos desidera- 2:26 For he longed after you all, and was full of heaviness,

because that ye had heard that he had been sick.

bat et maestus erat propterea quod audieratis illum infirma-

tum 2:27 nam et infirmatus est usque ad mortem sed Deus 2:27 For indeed he was sick nigh unto death: but God had mercy on him; and not on him only, but on me also, lest I should have sorrow upon sorrow.

misertus est eius non solum autem eius uerum etiam et mei ne tristitiam super tristitiam haberem 2:28 festinantius ergo 2:28 I sent him therefore the more carefully, that, when ye

see him again, ye may rejoice, and that I may be the less sorrowful.

misi illum ut uiso eo iterum gaudeatis et ego sine tristitia sim

2:29 excipite itaque illum cum omni gaudio in Domino et 2:29 Receive him therefore in the Lord with all gladness; and hold such in reputation:

Biblia Sacra Vulgata

2:30 Because for the work of Christ he was nigh unto eiusmodi cum honore habetote 2:30 quoniam propter opus

death, not regarding his life, to supply your lack of service toward me.

Christi usque ad mortem accessit tradens animam suam ut impleret id quod ex uobis deerat erga meum obsequium

3:1 Finally, my brethren, rejoice in the Lord. To write the 3:1 de cetero fratres mei gaudete in Domino eadem uobis same things to you, to me indeed is not grievous, but for you it is safe.

scribere mihi quidem non pigrum uobis autem necessarium

3:2 Beware of dogs, beware of evil workers, beware of the 3:2 uidete canes uidete malos operarios uidete concisionem

concision.

3:3 For we are the circumcision, which worship God in the 3:3 nos enim sumus circumcisio qui spiritu Deo seruimus et

spirit, and rejoice in Christ Jesus, and have no confidence in the flesh.

gloriamur in Christo Iesu et non in carne fiduciam habentes

3:4 Though I might also have confidence in the flesh. If 3:4 quamquam ego habeam confidentiam et in carne si quis

any other man thinketh that he hath whereof he might trust in the flesh, I more:

alius uidetur confidere in carne ego magis 3:5 circumcisus

3:5 Circumcised the eighth day, of the stock of Israel, of the tribe of Benjamin, an Hebrew of the Hebrews; as touching the law, a Pharisee;

octaua die ex genere Israhel de tribu Beniamin Hebraeus ex

3:6 Concerning zeal, persecuting the church; touching the Hebraeis secundum legem Pharisaeus 3:6 secundum aemurighteousness which is in the law, blameless.

lationem persequens ecclesiam Dei secundum iustitiam quae

3:7 But what things were gain to me, those I counted loss in lege est conuersatus sine querella 3:7 sed quae mihi fue-

for Christ.

runt lucra haec arbitratus sum propter Christum detrimenta

3:8 Yea doubtless, and I count all things but loss for the 3:8 uerumtamen existimo omnia detrimentum esse propter

excellency of the knowledge of Christ Jesus my Lord: for whom I have suffered the loss of all things, and do count

them but dung, that I may win Christ, eminentem scientiam Iesu Christi Domini mei propter quem

omnia detrimentum feci et arbitror ut stercora ut Christum

3:9 And be found in him, not having mine own righteous- lucri faciam 3:9 et inueniar in illo non habens meam iusti-

ness, which is of the law, but that which is through the faith of Christ, the righteousness which is of God by faith:

tiam quae ex lege est sed illam quae ex fide est Christi quae

3:10 That I may know him, and the power of his resur- ex Deo est iustitia in fide 3:10 ad agnoscendum illum et uir-

rection, and the fellowship of his sufferings, being made conformable unto his death;

tutem resurrectionis eius et societatem passionum illius con-

3:11 If by any means I might attain unto the resurrection figuratus morti eius 3:11 si quo modo occurram ad resurof the dead. 3:12 Not as though I had already attained, either were al- rectionem quae est ex mortuis 3:12 non quod iam accepeready perfect: but I follow after, if that I may apprehend that for which also I am apprehended of Christ Jesus.

rim aut iam perfectus sim sequor autem si conprehendam in

3:13 Brethren, I count not myself to have apprehended: quo et conprehensus sum a Christo Iesu 3:13 fratres ego me

but this one thing I do, forgetting those things which are behind, and reaching forth unto those things which are be-

fore, non arbitror conprehendisse unum autem quae quidem retro

sunt obliuiscens ad ea uero quae sunt in priora extendens me

3:14 I press toward the mark for the prize of the high call- 3:14 ad destinatum persequor ad brauium supernae uocatio-

ing of God in Christ Jesus.

3:15 Let us therefore, as many as be perfect, be thus nis Dei in Christo Iesu 3:15 quicumque ergo perfecti hoc

minded: and if in any thing ye be otherwise minded, God shall reveal even this unto you.

sentiamus et si quid aliter sapitis et hoc uobis Deus reue-

3:16 Nevertheless, whereto we have already attained, let labit 3:16 uerumtamen ad quod peruenimus ut idem sapiaus walk by the same rule, let us mind the same thing. 3:17 Brethren, be followers together of me, and mark them mus et in eadem permaneamus regula 3:17 imitatores mei which walk so as ye have us for an ensample.

Epistola Pauli Ad Philippenses

estote fratres et obseruate eos qui ita ambulant sicut habetis formam nos 3:18 multi enim ambulant quos saepe dice- 3:18 (For many walk, of whom I have told you often, and

now tell you even weeping, that they are the enemies of the cross of Christ:

bam uobis nunc autem et flens dico inimicos crucis Christi 3:19 quorum finis interitus quorum deus uenter et gloria in 3:19 Whose end is destruction, whose God is their belly,

and whose glory is in their shame, who mind earthly things.)

confusione ipsorum qui terrena sapiunt 3:20 nostra autem

3:20 For our conversation is in heaven; from whence also we look for the Saviour, the Lord Jesus Christ:

conuersatio in caelis est unde etiam saluatorem expectamus Dominum Iesum Christum 3:21 qui reformabit corpus hu- 3:21 Who shall change our vile body, that it may be fash-

ioned like unto his glorious body, according to the working whereby he is able even to subdue all things unto himself.

militatis nostrae configuratum corpori claritatis suae secundum operationem qua possit etiam subicere sibi omnia

4:1 itaque fratres mei carissimi et desiderantissimi gau- 4:1 Therefore, my brethren dearly beloved and longed for,

my joy and crown, so stand fast in the Lord, my dearly beloved.

dium meum et corona mea sic state in Domino carissimi

4:2 Euhodiam rogo et Syntychen deprecor id ipsum sapere 4:2 I beseech Euodias, and beseech Syntyche, that they be

of the same mind in the Lord.

in Domino 4:3 etiam rogo et te germane conpar adiuua il- 4:3 And I intreat thee also, true yokefellow, help those women which laboured with me in the gospel, with Clement also, and with other my fellowlabourers, whose

las quae mecum laborauerunt in euangelio cum Clemente et names are in the book of life. ceteris adiutoribus meis quorum nomina sunt in libro uitae 4:4 gaudete in Domino semper iterum dico gaudete 4:5 mo- 4:4 Rejoice in the Lord alway: and again I say, Rejoice.

4:5 Let your moderation be known unto all men. The Lord is at hand.

destia uestra nota sit omnibus hominibus Dominus prope 4:6 nihil solliciti sitis sed in omni oratione et obsecratione 4:6 Be careful for nothing; but in every thing by prayer and

supplication with thanksgiving let your requests be made known unto God.

cum gratiarum actione petitiones uestrae innotescant apud

Deum 4:7 et pax Dei quae exsuperat omnem sensum custo- 4:7 And the peace of God, which passeth all understand-

ing, shall keep your hearts and minds through Christ Jesus.

diat corda uestra et intellegentias uestras in Christo Iesu 4:8 4:8 Finally, brethren, whatsoever things are true, whatso-

ever things are honest, whatsoever things are just, whatsoever things are pure, whatsoever things are lovely, whatsoever things are of good report; if there be any virtue, and if there be any praise, think on these things.

de cetero fratres quaecumque sunt uera quaecumque pudica quaecumque iusta quaecumque sancta quaecumque amabilia quaecumque bonae famae si qua uirtus si qua laus haec cogitate 4:9 quae et didicistis et accepistis et audistis et uidis- 4:9 Those things, which ye have both learned, and re-

ceived, and heard, and seen in me, do: and the God of peace shall be with you.

tis in me haec agite et Deus pacis erit uobiscum 4:10 gaui-

4:10 But I rejoiced in the Lord greatly, that now at the last your care of me hath flourished again; wherein ye were also careful, but ye lacked opportunity.

sus sum autem in Domino uehementer quoniam tandem aliquando refloruistis pro me sentire sicut et sentiebatis occupati autem eratis 4:11 non quasi propter penuriam dico ego 4:11 Not that I speak in respect of want: for I have learned,

in whatsoever state I am, therewith to be content.

enim didici in quibus sum sufficiens esse 4:12 scio et hu- 4:12 I know both how to be abased, and I know how to

abound: every where and in all things I am instructed both to be full and to be hungry, both to abound and to suffer

miliari scio et abundare ubique et in omnibus institutus sum need.

et satiari et esurire et abundare et penuriam pati 4:13 om- 4:13 I can do all things through Christ which strengtheneth

me.

nia possum in eo qui me confortat 4:14 uerumtamen bene 4:14 Notwithstanding ye have well done, that ye did communicate with my affliction.

4:15 Now ye Philippians know also, that in the beginning fecistis communicantes tribulationi meae 4:15 scitis autem

of the gospel, when I departed from Macedonia, no church communicated with me as concerning giving and receiv-

ing, but ye only. et uos Philippenses quod in principio euangelii quando profectus sum a Macedonia nulla mihi ecclesia communicauit

4:16 For even in Thessalonica ye sent once and again unto in ratione dati et accepti nisi uos soli 4:16 quia et Thes-

my necessity. 4:17 Not because I desire a gift: but I desire fruit that may salonicam et semel et bis in usum mihi misistis 4:17 non abound to your account.

quia quaero datum sed requiro fructum abundantem in ra-

4:18 But I have all, and abound: I am full, having received tionem uestram 4:18 habeo autem omnia et abundo repleof Epaphroditus the things which were sent from you, an odour of a sweet smell, a sacrifice acceptable, wellpleas-

ing to God. tus sum acceptis ab Epafrodito quae misistis odorem sua-4:19 But my God shall supply all your need according to uitatis hostiam acceptam placentem Deo 4:19 Deus autem

his riches in glory by Christ Jesus.

meus impleat omne desiderium uestrum secundum diuitias

4:20 Now unto God and our Father be glory for ever and suas in gloria in Christo Iesu 4:20 Deo autem et Patri nos-

ever. Amen. 4:21 Salute every saint in Christ Jesus. The brethren which tro gloria in saecula saeculorum amen 4:21 salutate omnem are with me greet you.

sanctum in Christo Iesu salutant uos qui mecum sunt fratres

4:22 All the saints salute you, chiefly they that are of Cae- 4:22 salutant uos omnes sancti maxime autem qui de Caesasar’s household.

4:23 The grace of our Lord Jesus Christ be with you all. ris domo sunt 4:23 gratia Domini Iesu Christi cum spiritu

Amen.

uestro amen

Epistola Pauli Ad Colossenses

1:1 Paulus apostolus Christi Iesu per uoluntatem Dei et Ti- 1:1 Paul, an apostle of Jesus Christ by the will of God, and

Timotheus our brother,

motheus frater 1:2 his qui sunt Colossis sanctis et fidelibus 1:2 To the saints and faithful brethren in Christ which are

at Colosse: Grace be unto you, and peace, from God our Father and the Lord Jesus Christ.

fratribus in Christo Iesu gratia uobis et pax a Deo Patre nostro 1:3 gratias agimus Deo et Patri Domini nostri Iesu Ch- 1:3 We give thanks to God and the Father of our Lord Jesus

Christ, praying always for you,

risti semper pro uobis orantes 1:4 audientes fidem uestram 1:4 Since we heard of your faith in Christ Jesus, and of the

love which ye have to all the saints,

in Christo Iesu et dilectionem quam habetis in sanctos omnes 1:5 propter spem quae reposita est uobis in caelis quam au- 1:5 For the hope which is laid up for you in heaven,

whereof ye heard before in the word of the truth of the gospel;

distis in uerbo ueritatis euangelii 1:6 quod peruenit ad uos

1:6 Which is come unto you, as it is in all the world; and bringeth forth fruit, as it doth also in you, since the day ye heard of it, and knew the grace of God in truth:

sicut et in uniuerso mundo est et fructificat et crescit sicut in uobis ex ea die qua audistis et cognouistis gratiam Dei in ueritate 1:7 sicut didicistis ab Epaphra carissimo con- 1:7 As ye also learned of Epaphras our dear fellowservant,

who is for you a faithful minister of Christ;

seruo nostro qui est fidelis pro uobis minister Christi Iesu 1:8 qui etiam manifestauit nobis dilectionem uestram in Spi- 1:8 Who also declared unto us your love in the Spirit. ritu 1:9 ideo et nos ex qua die audiuimus non cessamus pro 1:9 For this cause we also, since the day we heard it, do

not cease to pray for you, and to desire that ye might be filled with the knowledge of his will in all wisdom and

uobis orantes et postulantes ut impleamini agnitione uolun- spiritual understanding; tatis eius in omni sapientia et intellectu spiritali 1:10 ut am- 1:10 That ye might walk worthy of the Lord unto all pleas-

ing, being fruitful in every good work, and increasing in the knowledge of God;

buletis digne Deo per omnia placentes in omni opere bono fructificantes et crescentes in scientia Dei 1:11 in omni uir- 1:11 Strengthened with all might, according to his glorious

power, unto all patience and longsuffering with joyfulness;

tute confortati secundum potentiam claritatis eius in omni patientia et longanimitate cum gaudio 1:12 gratias agentes 1:12 Giving thanks unto the Father, which hath made us

meet to be partakers of the inheritance of the saints in light:

Patri qui dignos nos fecit in partem sortis sanctorum in lumine 1:13 qui eripuit nos de potestate tenebrarum et trans- 1:13 Who hath delivered us from the power of darkness,

and hath translated us into the kingdom of his dear Son:

tulit in regnum Filii dilectionis suae 1:14 in quo habemus 1:14 In whom we have redemption through his blood, even

the forgiveness of sins:

redeeptionem remissionem peccatorum 1:15 qui est imago 1:15 Who is the image of the invisible God, the firstborn

of every creature:

Dei inuisibilis primogenitus omnis creaturae 1:16 quia in 1:16 For by him were all things created, that are in heaven,

and that are in earth, visible and invisible, whether they be thrones, or dominions, or principalities, or powers: all

ipso condita sunt uniuersa in caelis et in terra uisibilia et in- things were created by him, and for him: uisibilia siue throni siue dominationes siue principatus siue potestates omnia per ipsum et in ipso creata sunt 1:17 et 1:17 And he is before all things, and by him all things

consist.

ipse est ante omnes et omnia in ipso constant 1:18 et ipse 1:18 And he is the head of the body, the church: who is the

beginning, the firstborn from the dead; that in all things he might have the preeminence.

est caput corporis ecclesiae qui est principium primogenitus

ex mortuis ut sit in omnibus ipse primatum tenens 1:19 quia 1:19 For it pleased the Father that in him should all fulness

dwell;

in ipso conplacuit omnem plenitudinem habitare 1:20 et per 1:20 And, having made peace through the blood of his

cross, by him to reconcile all things unto himself; by him, I say, whether they be things in earth, or things in heaven.

eum reconciliare omnia in ipsum pacificans per sanguinem

Biblia Sacra Vulgata

1:21 And you, that were sometime alienated and enemies crucis eius siue quae in terris siue quae in caelis sunt 1:21

in your mind by wicked works, yet now hath he reconciled

et uos cum essetis aliquando alienati et inimici sensu in ope-

1:22 In the body of his flesh through death, to present you ribus malis 1:22 nunc autem reconciliauit in corpore carnis

holy and unblameable and unreproveable in his sight:

eius per mortem exhibere uos sanctos et inmaculatos et inre-

1:23 If ye continue in the faith grounded and settled, and prehensibiles coram ipso 1:23 si tamen permanetis in fide be not moved away from the hope of the gospel, which ye have heard, and which was preached to every creature which is under heaven; whereof I Paul am made a minister;

fundati et stabiles et inmobiles ab spe euangelii quod audistis quod praedicatum est in uniuersa creatura quae sub caelo est

1:24 Who now rejoice in my sufferings for you, and fill up cuius factus sum ego Paulus minister 1:24 qui nunc gaudeo

that which is behind of the afflictions of Christ in my flesh for his body’s sake, which is the church:

in passionibus pro uobis et adimpleo ea quae desunt passionum Christi in carne mea pro corpore eius quod est ecclesia

1:25 Whereof I am made a minister, according to the dis- 1:25 cuius factus sum ego minister secundum dispensationem

pensation of God which is given to me for you, to fulfil the word of God;

Dei quae data est mihi in uos ut impleam uerbum Dei 1:26

1:26 Even the mystery which hath been hid from ages and from generations, but now is made manifest to his saints:

mysterium quod absconditum fuit a saeculis et generationi-

1:27 To whom God would make known what is the riches bus nunc autem manifestatum est sanctis eius 1:27 quibus

of the glory of this mystery among the Gentiles; which is Christ in you, the hope of glory:

uoluit Deus notas facere diuitias gloriae sacramenti huius in

1:28 Whom we preach, warning every man, and teaching gentibus quod est Christus in uobis spes gloriae 1:28 quem

every man in all wisdom; that we may present every man perfect in Christ Jesus:

nos adnuntiamus corripientes omnem hominem et docentes omnem hominem in omni sapientia ut exhibeamus omnem

1:29 Whereunto I also labour, striving according to his hominem perfectum in Christo Iesu 1:29 in quo et laboro working, which worketh in me mightily.

certando secundum operationem eius quam operatur in me in uirtute

2:1 For I would that ye knew what great conflict I have for 2:1 uolo enim uos scire qualem sollicitudinem habeam pro

you, and for them at Laodicea, and for as many as have not seen my face in the flesh;

uobis et pro his qui sunt Laodiciae et quicumque non uide-

2:2 That their hearts might be comforted, being knit to- runt faciem meam in carne 2:2 ut consolentur corda ipsogether in love, and unto all riches of the full assurance of understanding, to the acknowledgement of the mystery of

God, and of the Father, and of Christ; rum instructi in caritate et in omnes diuitias plenitudinis in-2:3 In whom are hid all the treasures of wisdom and tellectus in agnitionem mysterii Dei Patris Christi Iesu 2:3

knowledge.

in quo sunt omnes thesauri sapientiae et scientiae abscon-

2:4 And this I say, lest any man should beguile you with diti 2:4 hoc autem dico ut nemo uos decipiat in subtilitate

enticing words. 2:5 For though I be absent in the flesh, yet am I with you in sermonum 2:5 nam et si corpore absens sum sed spiritu the spirit, joying and beholding your order, and the stedfastness of your faith in Christ.

uobiscum sum gaudens et uidens ordinem uestrum et fir-

2:6 As ye have therefore received Christ Jesus the Lord, mamentum eius quae in Christo est fidei uestrae 2:6 sicut so walk ye in him:

ergo accepistis Christum Iesum Dominum in ipso ambulate

2:7 Rooted and built up in him, and stablished in the faith, 2:7 radicati et superaedificati in ipso et confirmati fide sicut

as ye have been taught, abounding therein with thanksgiving.

Epistola Pauli Ad Colossenses

et didicistis abundantes in gratiarum actione 2:8 uidete ne 2:8 Beware lest any man spoil you through philosophy and

vain deceit, after the tradition of men, after the rudiments of the world, and not after Christ.

quis uos decipiat per philosophiam et inanem fallaciam secundum traditionem hominum secundum elementa mundi et non secundum Christum 2:9 quia in ipso inhabitat omnis 2:9 For in him dwelleth all the fulness of the Godhead

bodily.

plenitudo diuinitatis corporaliter 2:10 et estis in illo rep- 2:10 And ye are complete in him, which is the head of all

principality and power:

leti qui est caput omnis principatus et potestatis 2:11 in quo 2:11 In whom also ye are circumcised with the circumci-

sion made without hands, in putting off the body of the sins of the flesh by the circumcision of Christ:

et circumcisi estis circumcisione non manufacta in expoliatione corporis carnis in circumcisione Christi 2:12 conse- 2:12 Buried with him in baptism, wherein also ye are risen

with him through the faith of the operation of God, who hath raised him from the dead.

pulti ei in baptismo in quo et resurrexistis per fidem operationis Dei qui suscitauit illum a mortuis 2:13 et uos cum 2:13 And you, being dead in your sins and the uncircum-

cision of your flesh, hath he quickened together with him, having forgiven you all trespasses;

mortui essetis in delictis et praeputio carnis uestrae conuiuificauit cum illo donans uobis omnia delicta 2:14 delens 2:14 Blotting out the handwriting of ordinances that was

against us, which was contrary to us, and took it out of the way, nailing it to his cross;

quod aduersum nos erat chirografum decretis quod erat contrarium nobis et ipsum tulit de medio adfigens illud cruci 2:15 expolians principatus et potestates traduxit palam trium- 2:15 And having spoiled principalities and powers, he

made a shew of them openly, triumphing over them in it.

phans illos in semet ipso 2:16 nemo ergo uos iudicet in cibo 2:16 Let no man therefore judge you in meat, or in drink,

or in respect of an holyday, or of the new moon, or of the sabbath days:

aut in potu aut in parte diei festi aut neomeniae aut sabbatorum 2:17 quae sunt umbra futurorum corpus autem Christi 2:17 Which are a shadow of things to come; but the body

is of Christ.

2:18 nemo uos seducat uolens in humilitate et religione an- 2:18 Let no man beguile you of your reward in a voluntary

humility and worshipping of angels, intruding into those things which he hath not seen, vainly puffed up by his

gelorum quae non uidit ambulans frustra inflatus sensu car- fleshly mind, nis suae 2:19 et non tenens caput ex quo totum corpus per 2:19 And not holding the Head, from which all the body

by joints and bands having nourishment ministered, and knit together, increaseth with the increase of God.

nexus et coniunctiones subministratum et constructum crescit in augmentum Dei 2:20 si mortui estis cum Christo ab 2:20 Wherefore if ye be dead with Christ from the rudi-

ments of the world, why, as though living in the world, are ye subject to ordinances,

elementis mundi quid adhuc tamquam uiuentes in mundo decernitis 2:21 ne tetigeris neque gustaueris neque contrec- 2:21 (Touch not; taste not; handle not; taueris 2:22 quae sunt omnia in interitu ipso usu secundum 2:22 Which all are to perish with the using;) after the com-

mandments and doctrines of men?

praecepta et doctrinas hominum 2:23 quae sunt rationem 2:23 Which things have indeed a shew of wisdom in will

worship, and humility, and neglecting of the body: not in any honour to the satisfying of the flesh.

quidem habentia sapientiae in superstitione et humilitate et ad non parcendum corpori non in honore aliquo ad saturitatem carnis

3:1 igitur si conresurrexistis Christo quae sursum sunt 3:1 If ye then be risen with Christ, seek those things which

are above, where Christ sitteth on the right hand of God.

quaerite ubi Christus est in dextera Dei sedens 3:2 quae 3:2 Set your affection on things above, not on things on

the earth.

sursum sunt sapite non quae supra terram 3:3 mortui enim 3:3 For ye are dead, and your life is hid with Christ in God.

Biblia Sacra Vulgata

3:4 When Christ, who is our life, shall appear, then shall estis et uita uestra abscondita est cum Christo in Deo 3:4 ye also appear with him in glory.

cum Christus apparuerit uita uestra tunc et uos apparebitis

3:5 Mortify therefore your members which are upon the cum ipso in gloria 3:5 mortificate ergo membra uestra quae

earth; fornication, uncleanness, inordinate affection, evil concupiscence, and covetousness, which is idolatry:

sunt super terram fornicationem inmunditiam libidinem concupiscentiam malam et auaritiam quae est simulacrorum ser-

3:6 For which things’ sake the wrath of God cometh on uitus 3:6 propter quae uenit ira Dei super filios incredulita-

the children of disobedience:

3:7 In the which ye also walked some time, when ye lived tis 3:7 in quibus et uos ambulastis aliquando cum uiueretis

in them. 3:8 But now ye also put off all these; anger, wrath, malice, in illis 3:8 nunc autem deponite et uos omnia iram indigblasphemy, filthy communication out of your mouth.

nationem malitiam blasphemiam turpem sermonem de ore

3:9 Lie not one to another, seeing that ye have put off the uestro 3:9 nolite mentiri inuicem expoliantes uos ueterem

old man with his deeds;

3:10 And have put on the new man, which is renewed in hominem cum actibus eius 3:10 et induentes nouum eum

knowledge after the image of him that created him:

qui renouatur in agnitionem secundum imaginem eius qui

3:11 Where there is neither Greek nor Jew, circumcision creauit eum 3:11 ubi non est gentilis et Iudaeus circumcisio

nor uncircumcision, Barbarian, Scythian, bond nor free: but Christ is all, and in all.

et praeputium barbarus et Scytha seruus et liber sed omnia et

3:12 Put on therefore, as the elect of God, holy and in omnibus Christus 3:12 induite uos ergo sicut electi Dei

beloved, bowels of mercies, kindness, humbleness of mind, meekness, longsuffering;

sancti et dilecti uiscera misericordiae benignitatem humili-

3:13 Forbearing one another, and forgiving one another, if tatem modestiam patientiam 3:13 subportantes inuicem et any man have a quarrel against any: even as Christ forgave you, so also do ye.

donantes uobis ipsis si quis aduersus aliquem habet querel-

3:14 And above all these things put on charity, which is lam sicut et Dominus donauit uobis ita et uos 3:14 super the bond of perfectness.

omnia autem haec caritatem quod est uinculum perfectionis

3:15 And let the peace of God rule in your hearts, to the 3:15 et pax Christi exultet in cordibus uestris in qua et uocati

which also ye are called in one body; and be ye thankful.

3:16 Let the word of Christ dwell in you richly in all wis- estis in uno corpore et grati estote 3:16 uerbum Christi ha-

dom; teaching and admonishing one another in psalms and hymns and spiritual songs, singing with grace in your

hearts to the Lord. bitet in uobis abundanter in omni sapientia docentes et com-

monentes uosmet ipsos psalmis hymnis canticis spiritalibus

3:17 And whatsoever ye do in word or deed, do all in the in gratia cantantes in cordibus uestris Deo 3:17 omne quod-

name of the Lord Jesus, giving thanks to God and the Father by him.

cumque facitis in uerbo aut in opere omnia in nomine Do-

3:18 Wives, submit yourselves unto your own husbands, mini Iesu gratias agentes Deo et Patri per ipsum 3:18 mu-

as it is fit in the Lord.

3:19 Husbands, love your wives, and be not bitter against lieres subditae estote uiris sicut oportet in Domino 3:19 uiri

them.

3:20 Children, obey your parents in all things: for this is diligite uxores et nolite amari esse ad illas 3:20 filii oboe-

well pleasing unto the Lord.

dite parentibus per omnia hoc enim placitum est in Domino

3:21 Fathers, provoke not your children to anger, lest they 3:21 patres nolite ad indignationem prouocare filios uestros

be discouraged. 3:22 Servants, obey in all things your masters according ut non pusillo animo fiant 3:22 serui oboedite per omnia to the flesh; not with eyeservice, as menpleasers; but in singleness of heart, fearing God;

dominis carnalibus non ad oculum seruientes quasi homini-

Epistola Pauli Ad Colossenses

bus placentes sed in simplicitate cordis timentes Dominum 3:23 quodcumque facitis ex animo operamini sicut Domino 3:23 And whatsoever ye do, do it heartily, as to the Lord,

and not unto men;

et non hominibus 3:24 scientes quod a Domino accipietis 3:24 Knowing that of the Lord ye shall receive the reward

of the inheritance: for ye serve the Lord Christ.

retributionem hereditatis Domino Christo seruite 3:25 qui 3:25 But he that doeth wrong shall receive for the wrong

which he hath done: and there is no respect of persons.

enim iniuriam facit recipiet id quod inique gessit et non est personarum acceptio

4:1 domini quod iustum est et aequum seruis praestate 4:1 Masters, give unto your servants that which is just and

equal; knowing that ye also have a Master in heaven.

scientes quoniam et uos Dominum habetis in caelo 4:2 ora- 4:2 Continue in prayer, and watch in the same with thanks-

giving;

tioni instate uigilantes in ea in gratiarum actione 4:3 oran- 4:3 Withal praying also for us, that God would open unto

us a door of utterance, to speak the mystery of Christ, for which I am also in bonds:

tes simul et pro nobis ut Deus aperiat nobis ostium sermonis ad loquendum mysterium Christi propter quod etiam uinctus sum 4:4 ut manifestem illud ita ut oportet me loqui 4:5 in 4:4 That I may make it manifest, as I ought to speak.

4:5 Walk in wisdom toward them that are without, redeeming the time.

sapientia ambulate ad eos qui foris sunt tempus rediientes 4:6 sermo uester semper in gratia sale sit conditus ut scia- 4:6 Let your speech be alway with grace, seasoned with

salt, that ye may know how ye ought to answer every man.

tis quomodo oporteat uos unicuique respondere 4:7 quae 4:7 All my state shall Tychicus declare unto you, who is a

beloved brother, and a faithful minister and fellowservant in the Lord:

circa me sunt omnia uobis nota faciet Tychicus carissimus frater et fidelis minister et conseruus in Domino 4:8 quem 4:8 Whom I have sent unto you for the same purpose, that

he might know your estate, and comfort your hearts;

misi ad uos ad hoc ipsum ut cognoscat quae circa uos sunt

et consoletur corda uestra 4:9 cum Onesimo carissimo et 4:9 With Onesimus, a faithful and beloved brother, who is one of you. They shall make known unto you all things which are done here.

fideli fratre qui est ex uobis omnia quae hic aguntur nota facient uobis 4:10 salutat uos Aristarchus concaptiuus meus 4:10 Aristarchus my fellowprisoner saluteth you, and Mar-

cus, sister’s son to Barnabas, (touching whom ye received commandments: if he come unto you, receive him;)

et Marcus consobrinus Barnabae de quo accepistis mandata si uenerit ad uos excipite illum 4:11 et Iesus qui dicitur Ius- 4:11 And Jesus, which is called Justus, who are of the

circumcision. These only are my fellowworkers unto the kingdom of God, which have been a comfort unto me.

tus qui sunt ex circumcisione hii soli sunt adiutores in regno Dei qui mihi fuerunt solacio 4:12 salutat uos Epaphras qui 4:12 Epaphras, who is one of you, a servant of Christ,

saluteth you, always labouring fervently for you in prayers, that ye may stand perfect and complete in all the

ex uobis est seruus Christi Iesu semper sollicitus pro uobis will of God. in orationibus ut stetis perfecti et pleni in omni uoluntate Dei 4:13 testimonium enim illi perhibeo quod habet multum labo- 4:13 For I bear him record, that he hath a great zeal for

you, and them that are in Laodicea, and them in Hierapolis.

rem pro uobis et pro his qui sunt Laodiciae et qui Hierapoli 4:14 salutat uos Lucas medicus carissimus et Demas 4:15 4:14 Luke, the beloved physician, and Demas, greet you.

4:15 Salute the brethren which are in Laodicea, and Nymphas, and the church which is in his house.

salutate fratres qui sunt Laodiciae et Nympham et quae in domo eius est ecclesiam 4:16 et cum lecta fuerit apud uos 4:16 And when this epistle is read among you, cause that

it be read also in the church of the Laodiceans; and that ye likewise read the epistle from Laodicea.

epistula facite ut et in Laodicensium ecclesia legatur et eam 4:17 And say to Archippus, Take heed to the ministry quae Laodicensium est uos legatis 4:17 et dicite Archippo

which thou hast received in the Lord, that thou fulfil it.

uide ministerium quod accepisti in Domino ut illud impleas

4:18 The salutation by the hand of me Paul. Remember 4:18 salutatio mea manu Pauli memores estote uinculorum

my bonds. Grace be with you. Amen.

meorum gratia uobiscum amen

Epistola Pauli Ad Thessalonicenses I

1:1 Paulus et Siluanus et Timotheus ecclesiae Thessaloni- 1:1 Paul, and Silvanus, and Timotheus, unto the church of

the Thessalonians which is in God the Father and in the Lord Jesus Christ: Grace be unto you, and peace, from

censium in Deo Patre et Domino Iesu Christo gratia uobis et God our Father, and the Lord Jesus Christ. pax 1:2 gratias agimus Deo semper pro omnibus uobis me- 1:2 We give thanks to God always for you all, making

mention of you in our prayers;

moriam facientes in orationibus nostris sine intermissione 1:3 memores operis fidei uestrae et laboris et caritatis et su- 1:3 Remembering without ceasing your work of faith, and

labour of love, and patience of hope in our Lord Jesus Christ, in the sight of God and our Father;

stinentiae spei Domini nostri Iesu Christi ante Deum et Patrem nostrum 1:4 scientes fratres dilecti a Deo electionem 1:4 Knowing, brethren beloved, your election of God. uestram 1:5 quia euangelium nostrum non fuit ad uos in 1:5 For our gospel came not unto you in word only, but

also in power, and in the Holy Ghost, and in much assurance; as ye know what manner of men we were among

sermone tantum sed et in uirtute et in Spiritu Sancto et in ple- you for your sake. nitudine multa sicut scitis quales fuerimus uobis propter uos 1:6 et uos imitatores nostri facti estis et Domini excipien- 1:6 And ye became followers of us, and of the Lord, hav-

ing received the word in much affliction, with joy of the Holy Ghost.

tes uerbum in tribulatione multa cum gaudio Spiritus Sancti 1:7 ita ut facti sitis forma omnibus credentibus in Macedonia 1:7 So that ye were ensamples to all that believe in Mace-

donia and Achaia.

et in Achaia 1:8 a uobis enim diffamatus est sermo Domini 1:8 For from you sounded out the word of the Lord not

only in Macedonia and Achaia, but also in every place your faith to God-ward is spread abroad; so that we need

non solum in Macedonia et in Achaia sed in omni loco fi- not to speak any thing. des uestra quae est ad Deum profecta est ita ut non sit nobis necesse quicquam loqui 1:9 ipsi enim de nobis adnuntiant 1:9 For they themselves shew of us what manner of enter-

ing in we had unto you, and how ye turned to God from idols to serve the living and true God;

qualem introitum habuerimus ad uos et quomodo conuersi estis ad Deum a simulacris seruire Deo uiuo et uero 1:10 et 1:10 And to wait for his Son from heaven, whom he raised

from the dead, even Jesus, which delivered us from the wrath to come.

expectare Filium eius de caelis quem suscitauit ex mortuis Iesum qui eripuit nos ab ira uentura

2:1 nam ipsi scitis fratres introitum nostrum ad uos quia 2:1 For yourselves, brethren, know our entrance in unto

you, that it was not in vain:

non inanis fuit 2:2 sed ante passi et contumeliis affecti sic- 2:2 But even after that we had suffered before, and were

shamefully entreated, as ye know, at Philippi, we were bold in our God to speak unto you the gospel of God with

ut scitis in Philippis fiduciam habuimus in Deo nostro loqui much contention. ad uos euangelium Dei in multa sollicitudine 2:3 exhortatio 2:3 For our exhortation was not of deceit, nor of unclean-

ness, nor in guile:

enim nostra non de errore neque de inmunditia neque in dolo 2:4 sed sicut probati sumus a Deo ut crederetur nobis euan- 2:4 But as we were allowed of God to be put in trust with

the gospel, even so we speak; not as pleasing men, but God, which trieth our hearts.

gelium ita loquimur non quasi hominibus placentes sed Deo qui probat corda nostra 2:5 neque enim aliquando fuimus 2:5 For neither at any time used we flattering words, as ye

know, nor a cloke of covetousness; God is witness:

in sermone adulationis sicut scitis neque in occasione auaritiae Deus testis est 2:6 nec quaerentes ab hominibus glo- 2:6 Nor of men sought we glory, neither of you, nor yet

of others, when we might have been burdensome, as the apostles of Christ.

riam neque a uobis neque ab aliis 2:7 cum possimus oneri

2:7 But we were gentle among you, even as a nurse cherisheth her children:

esse ut Christi apostoli sed facti sumus lenes in medio ue-

Biblia Sacra Vulgata

2:8 So being affectionately desirous of you, we were will- strum tamquam si nutrix foueat filios suos 2:8 ita desideing to have imparted unto you, not the gospel of God only, but also our own souls, because ye were dear unto us.

rantes uos cupide uolebamus tradere uobis non solum euangelium Dei sed etiam animas nostras quoniam carissimi no-

2:9 For ye remember, brethren, our labour and travail: for bis facti estis 2:9 memores enim estis fratres laborem nolabouring night and day, because we would not be chargeable unto any of you, we preached unto you the gospel of

God. strum et fatigationem nocte et die operantes ne quem ue-

strum grauaremus praedicauimus in uobis euangelium Dei

2:10 Ye are witnesses, and God also, how holily and justly 2:10 uos testes estis et Deus quam sancte et iuste et sine que-

and unblameably we behaved ourselves among you that believe:

rella uobis qui credidistis fuimus 2:11 sicut scitis qualiter

2:11 As ye know how we exhorted and comforted and charged every one of you, as a father doth his children,

2:12 That ye would walk worthy of God, who hath called unumquemque uestrum tamquam pater filios suos 2:12 de-

you unto his kingdom and glory.

precantes uos et consolantes testificati sumus ut ambularetis

2:13 For this cause also thank we God without ceasing, be- digne Deo qui uocauit uos in suum regnum et gloriam 2:13 cause, when ye received the word of God which ye heard of us, ye received it not as the word of men, but as it is in truth, the word of God, which effectually worketh also in you that believe.

ideo et nos gratias agimus Deo sine intermissione quoniam cum accepissetis a nobis uerbum auditus Dei accepistis non ut uerbum hominum sed sicut est uere uerbum Dei qui ope-

2:14 For ye, brethren, became followers of the churches of ratur in uobis qui credidistis 2:14 uos enim imitatores facti

God which in Judaea are in Christ Jesus: for ye also have suffered like things of your own countrymen, even as they

have of the Jews: estis fratres ecclesiarum Dei quae sunt in Iudaea in Christo

Iesu quia eadem passi estis et uos a contribulibus uestris sic-

2:15 Who both killed the Lord Jesus, and their own ut et ipsi a Iudaeis 2:15 qui et Dominum occiderunt Iesum

prophets, and have persecuted us; and they please not God, and are contrary to all men:

et prophetas et nos persecuti sunt et Deo non placent et om-

2:16 Forbidding us to speak to the Gentiles that they might nibus hominibus aduersantur 2:16 prohibentes nos gentibus

be saved, to fill up their sins alway: for the wrath is come upon them to the uttermost.

loqui ut saluae fiant ut impleant peccata sua semper praeue-

2:17 But we, brethren, being taken from you for a short nit autem ira Dei super illos usque in finem 2:17 nos autem

time in presence, not in heart, endeavoured the more abundantly to see your face with great desire.

fratres desolati a uobis ad tempus horae aspectu non corde abundantius festinauimus faciem uestram uidere cum multo

2:18 Wherefore we would have come unto you, even I desiderio 2:18 quoniam uoluimus uenire ad uos ego qui-

Paul, once and again; but Satan hindered us.

2:19 For what is our hope, or joy, or crown of rejoicing? dem Paulus et semel et iterum et inpediuit nos Satanas 2:19

Are not even ye in the presence of our Lord Jesus Christ at his coming?

quae est enim nostra spes aut gaudium aut corona gloriae nonne uos ante Dominum nostrum Iesum estis in aduentu

2:20 For ye are our glory and joy. eius 2:20 uos enim estis gloria nostra et gaudium

3:1 Wherefore when we could no longer forbear, we 3:1 propter quod non sustinentes amplius placuit nobis re-

thought it good to be left at Athens alone; 3:2 And sent Timotheus, our brother, and minister of God, manere Athenis solis 3:2 et misimus Timotheum fratrem and our fellowlabourer in the gospel of Christ, to establish you, and to comfort you concerning your faith:

nostrum et ministrum Dei in euangelio Christi ad confirman-

3:3 That no man should be moved by these afflictions: for dos uos et exhortandos pro fide uestra 3:3 ut nemo mouea-

yourselves know that we are appointed thereunto.

Epistola Pauli Ad Thessalonicenses I

tur in tribulationibus istis ipsi enim scitis quod in hoc positi

sumus 3:4 nam et cum apud uos essemus praedicebamus 3:4 For verily, when we were with you, we told you before that we should suffer tribulation; even as it came to pass, and ye know.

uobis passuros nos tribulationes sicut et factum est et scitis 3:5 propterea et ego amplius non sustinens misi ad co- 3:5 For this cause, when I could no longer forbear, I sent

to know your faith, lest by some means the tempter have tempted you, and our labour be in vain.

gnoscendam fidem uestram ne forte temptauerit uos is qui temptat et inanis fiat labor noster 3:6 nunc autem ueniente 3:6 But now when Timotheus came from you unto us, and

brought us good tidings of your faith and charity, and that ye have good remembrance of us always, desiring greatly

Timotheo ad nos a uobis et adnuntiante nobis fidem et carita- to see us, as we also to see you: tem uestram et quia memoriam nostri habetis bonam semper desiderantes nos uidere sicut nos quoque uos 3:7 ideo con- 3:7 Therefore, brethren, we were comforted over you in

all our affliction and distress by your faith:

solati sumus fratres in uobis in omni necessitate et tribulatione nostra per uestram fidem 3:8 quoniam nunc uiuimus 3:8 For now we live, if ye stand fast in the Lord. si uos statis in Domino 3:9 quam enim gratiarum actionem 3:9 For what thanks can we render to God again for you,

for all the joy wherewith we joy for your sakes before our God;

possumus Deo retribuere pro uobis in omni gaudio quo gau-

demus propter uos ante Deum nostrum 3:10 nocte et die 3:10 Night and day praying exceedingly that we might see your face, and might perfect that which is lacking in your faith?

abundantius orantes ut uideamus faciem uestram et conpleamus ea quae desunt fidei uestrae 3:11 ipse autem Deus et 3:11 Now God himself and our Father, and our Lord Jesus

Christ, direct our way unto you.

Pater noster et Dominus Iesus dirigat uiam nostram ad uos 3:12 uos autem Dominus multiplicet et abundare faciat cari- 3:12 And the Lord make you to increase and abound in

love one toward another, and toward all men, even as we do toward you:

tatem in inuicem et in omnes quemadmodum et nos in uobis 3:13 ad confirmanda corda uestra sine querella in sanctitate 3:13 To the end he may stablish your hearts unblameable

in holiness before God, even our Father, at the coming of our Lord Jesus Christ with all his saints.

ante Deum et Patrem nostrum in aduentu Domini nostri Iesu cum omnibus sanctis eius amen

4:1 de cetero ergo fratres rogamus uos et obsecramus in 4:1 Furthermore then we beseech you, brethren, and ex-

hort you by the Lord Jesus, that as ye have received of us how ye ought to walk and to please God, so ye would

Domino Iesu ut quemadmodum accepistis a nobis quomodo abound more and more. uos oporteat ambulare et placere Deo sicut et ambulatis ut abundetis magis 4:2 scitis enim quae praecepta dederimus 4:2 For ye know what commandments we gave you by the

Lord Jesus.

uobis per Dominum Iesum 4:3 haec est enim uoluntas Dei 4:3 For this is the will of God, even your sanctification,

that ye should abstain from fornication:

sanctificatio uestra 4:4 ut abstineatis uos a fornicatione ut 4:4 That every one of you should know how to possess his

vessel in sanctification and honour;

sciat unusquisque uestrum suum uas possidere in sanctificatione et honore 4:5 non in passione desiderii sicut et gentes 4:5 Not in the lust of concupiscence, even as the Gentiles

which know not God:

quae ignorant Deum 4:6 ut ne quis supergrediatur neque 4:6 That no man go beyond and defraud his brother in any

matter: because that the Lord is the avenger of all such, as we also have forewarned you and testified.

circumueniat in negotio fratrem suum quoniam uindex est Dominus de his omnibus sicut et praediximus uobis et testi-

Biblia Sacra Vulgata

4:7 For God hath not called us unto uncleanness, but unto ficati sumus 4:7 non enim uocauit nos Deus in inmunditia holiness.

4:8 He therefore that despiseth, despiseth not man, but sed in sanctificatione 4:8 itaque qui spernit non hominem

God, who hath also given unto us his holy Spirit.

spernit sed Deum qui etiam dedit Spiritum suum Sanctum

4:9 But as touching brotherly love ye need not that I write in uobis 4:9 de caritate autem fraternitatis non necesse ha-

unto you: for ye yourselves are taught of God to love one another.

bemus scribere uobis ipsi enim uos a Deo didicistis ut dili-

4:10 And indeed ye do it toward all the brethren which are gatis inuicem 4:10 etenim facitis illud in omnes fratres in

in all Macedonia: but we beseech you, brethren, that ye increase more and more;

uniuersa Macedonia rogamus autem uos fratres ut abunde-

4:11 And that ye study to be quiet, and to do your own tis magis 4:11 et operam detis ut quieti sitis et ut uestrum

business, and to work with your own hands, as we commanded you;

negotium agatis et operemini manibus uestris sicut praece-

4:12 That ye may walk honestly toward them that are with- pimus uobis 4:12 et ut honeste ambuletis ad eos qui foris out, and that ye may have lack of nothing.

4:13 But I would not have you to be ignorant, brethren, sunt et nullius aliquid desideretis 4:13 nolumus autem uos

concerning them which are asleep, that ye sorrow not, even as others which have no hope.

ignorare fratres de dormientibus ut non contristemini sicut

4:14 For if we believe that Jesus died and rose again, even et ceteri qui spem non habent 4:14 si enim credimus quod so them also which sleep in Jesus will God bring with him.

Iesus mortuus est et resurrexit ita et Deus eos qui dormierunt

4:15 For this we say unto you by the word of the Lord, per Iesum adducet cum eo 4:15 hoc enim uobis dicimus in

that we which are alive and remain unto the coming of the Lord shall not prevent them which are asleep.

uerbo Domini quia nos qui uiuimus qui residui sumus in ad-

4:16 For the Lord himself shall descend from heaven with uentum Domini non praeueniemus eos qui dormierunt 4:16 a shout, with the voice of the archangel, and with the trump of God: and the dead in Christ shall rise first:

quoniam ipse Dominus in iussu et in uoce archangeli et in tuba Dei descendet de caelo et mortui qui in Christo sunt re-

4:17 Then we which are alive and remain shall be caught surgent primi 4:17 deinde nos qui uiuimus qui relinquimur

up together with them in the clouds, to meet the Lord in the air: and so shall we ever be with the Lord.

simul rapiemur cum illis in nubibus obuiam Domino in aera

4:18 Wherefore comfort one another with these words. et sic semper cum Domino erimus 4:18 itaque consolamini inuicem in uerbis istis

5:1 But of the times and the seasons, brethren, ye have no 5:1 de temporibus autem et momentis fratres non indigeneed that I write unto you.

5:2 For yourselves know perfectly that the day of the Lord tis ut scribamus uobis 5:2 ipsi enim diligenter scitis quia

so cometh as a thief in the night.

5:3 For when they shall say, Peace and safety; then sudden dies Domini sicut fur in nocte ita ueniet 5:3 cum enim dixe-

destruction cometh upon them, as travail upon a woman with child; and they shall not escape.

rint pax et securitas tunc repentinus eis superueniet interi-

5:4 But ye, brethren, are not in darkness, that that day tus sicut dolor in utero habenti et non effugient 5:4 uos au-

should overtake you as a thief.

tem fratres non estis in tenebris ut uos dies ille tamquam

5:5 Ye are all the children of light, and the children of the fur conprehendat 5:5 omnes enim uos filii lucis estis et fiday: we are not of the night, nor of darkness. 5:6 Therefore let us not sleep, as do others; but let us watch lii diei non sumus noctis neque tenebrarum 5:6 igitur non and be sober. 5:7 For they that sleep sleep in the night; and they that be dormiamus sicut ceteri sed uigilemus et sobrii simus 5:7 drunken are drunken in the night.

qui enim dormiunt nocte dormiunt et qui ebrii sunt nocte eb-

Epistola Pauli Ad Thessalonicenses I

rii sunt 5:8 nos autem qui diei sumus sobrii simus induti 5:8 But let us, who are of the day, be sober, putting on the

breastplate of faith and love; and for an helmet, the hope of salvation.

loricam fidei et caritatis et galeam spem salutis 5:9 quo-

5:9 For God hath not appointed us to wrath, but to obtain salvation by our Lord Jesus Christ,

niam non posuit nos Deus in iram sed in adquisitionem salutis per Dominum nostrum Iesum Christum 5:10 qui mor- 5:10 Who died for us, that, whether we wake or sleep, we

should live together with him.

tuus est pro nobis ut siue uigilemus siue dormiamus simul cum illo uiuamus 5:11 propter quod consolamini inuicem 5:11 Wherefore comfort yourselves together, and edify

one another, even as also ye do.

et aedificate alterutrum sicut et facitis 5:12 rogamus autem 5:12 And we beseech you, brethren, to know them which

labour among you, and are over you in the Lord, and admonish you;

uos fratres ut noueritis eos qui laborant inter uos et praesunt uobis in Domino et monent uos 5:13 ut habeatis il- 5:13 And to esteem them very highly in love for their

work’s sake. And be at peace among yourselves.

los abundantius in caritate propter opus illorum pacem habete cum eis 5:14 rogamus autem uos fratres corripite in- 5:14 Now we exhort you, brethren, warn them that are un-

ruly, comfort the feebleminded, support the weak, be patient toward all men.

quietos consolamini pusillianimes suscipite infirmos patien-

tes estote ad omnes 5:15 uidete ne quis malum pro malo 5:15 See that none render evil for evil unto any man; but ever follow that which is good, both among yourselves, and to all men.

alicui reddat sed semper quod bonum est sectamini et in inuicem et in omnes 5:16 semper gaudete 5:17 sine intermis- 5:16 Rejoice evermore.

5:17 Pray without ceasing.

sione orate 5:18 in omnibus gratias agite haec enim uolun- _{5:18 In every thing give thanks: for this is the will of God}

in Christ Jesus concerning you.

tas Dei est in Christo Iesu in omnibus uobis 5:19 Spiritum 5:19 Quench not the Spirit. nolite extinguere 5:20 prophetias nolite spernere 5:21 om- 5:20 Despise not prophesyings.

5:21 Prove all things; hold fast that which is good.

nia autem probate quod bonum est tenete 5:22 ab omni spe- _{5:22 Abstain from all appearance of evil.} cie mala abstinete uos 5:23 ipse autem Deus pacis sanctifi- 5:23 And the very God of peace sanctify you wholly; and I

pray God your whole spirit and soul and body be preserved blameless unto the coming of our Lord Jesus Christ.

cet uos per omnia et integer spiritus uester et anima et corpus sine querella in aduentu Domini nostri Iesu Christi seruetur 5:24 fidelis est qui uocauit uos qui etiam faciet 5:25 fratres 5:24 Faithful is he that calleth you, who also will do it.

5:25 Brethren, pray for us.

orate pro nobis 5:26 salutate fratres omnes in osculo sancto _{5:26 Greet all the brethren with an holy kiss.}

5:27 adiuro uos per Dominum ut legatur epistula omnibus 5:27 I charge you by the Lord that this epistle be read unto

all the holy brethren.

sanctis fratribus 5:28 gratia Domini nostri Iesu Christi uo- 5:28 The grace of our Lord Jesus Christ be with you.

Amen.

biscum amen

Biblia Sacra Vulgata

MDCCIV

Epistola Pauli Ad Thessalonicenses II

1:1 Paulus et Siluanus et Timotheus ecclesiae Thessalo- 1:1 Paul, and Silvanus, and Timotheus, unto the church of

the Thessalonians in God our Father and the Lord Jesus Christ:

nicensium in Deo Patre nostro et Domino Iesu Christo 1:2

1:2 Grace unto you, and peace, from God our Father and the Lord Jesus Christ.

gratia uobis et pax a Deo Patre nostro et Domino Iesu Christo 1:3 gratias agere debemus Deo semper pro uobis fra- 1:3 We are bound to thank God always for you, brethren,

as it is meet, because that your faith groweth exceedingly, and the charity of every one of you all toward each other

tres ita ut dignum est quoniam supercrescit fides uestra et aboundeth; abundat caritas uniuscuiusque omnium uestrum in inuicem 1:4 ita ut et nos ipsi in uobis gloriemur in ecclesiis Dei pro 1:4 So that we ourselves glory in you in the churches of

God for your patience and faith in all your persecutions and tribulations that ye endure:

patientia uestra et fide in omnibus persecutionibus uestris et tribulationibus quas sustinetis 1:5 in exemplum iusti iudi- 1:5 Which is a manifest token of the righteous judgment

of God, that ye may be counted worthy of the kingdom of God, for which ye also suffer:

cii Dei ut digni habeamini regno Dei pro quo et patimini 1:6 si tamen iustum est apud Deum retribuere tribulationem 1:6 Seeing it is a righteous thing with God to recompense

tribulation to them that trouble you;

his qui uos tribulant 1:7 et uobis qui tribulamini requiem 1:7 And to you who are troubled rest with us, when the

Lord Jesus shall be revealed from heaven with his mighty angels,

nobiscum in reuelatione Domini Iesu de caelo cum angelis uirtutis eius 1:8 in flamma ignis dantis uindictam his qui 1:8 In flaming fire taking vengeance on them that know

not God, and that obey not the gospel of our Lord Jesus Christ:

non nouerunt Deum et qui non oboediunt euangelio Domini nostri Iesu 1:9 qui poenas dabunt in interitu aeternas a facie 1:9 Who shall be punished with everlasting destruction

from the presence of the Lord, and from the glory of his power;

Domini et a gloria uirtutis eius 1:10 cum uenerit glorificari

1:10 When he shall come to be glorified in his saints, and to be admired in all them that believe (because our testimony among you was believed) in that day.

in sanctis suis et admirabilis fieri in omnibus qui crediderunt quia creditum est testimonium nostrum super uos in die illo 1:11 in quo etiam oramus semper pro uobis ut dignetur uos 1:11 Wherefore also we pray always for you, that our God

would count you worthy of this calling, and fulfil all the good pleasure of his goodness, and the work of faith with

uocatione sua Deus et impleat omnem uoluntatem bonitatis power: et opus fidei in uirtute 1:12 ut clarificetur nomen Domini 1:12 That the name of our Lord Jesus Christ may be glo-

rified in you, and ye in him, according to the grace of our God and the Lord Jesus Christ.

nostri Iesu Christi in uobis et uos in illo secundum gratiam Dei nostri et Domini Iesu Christi

2:1 rogamus autem uos fratres per aduentum Domini nos- 2:1 Now we beseech you, brethren, by the coming of our

Lord Jesus Christ, and by our gathering together unto him,

tri Iesu Christi et nostrae congregationis in ipsum 2:2 ut 2:2 That ye be not soon shaken in mind, or be troubled,

neither by spirit, nor by word, nor by letter as from us, as that the day of Christ is at hand.

non cito moueamini a sensu neque terreamini neque per spiritum neque per sermonem neque per epistulam tamquam per nos quasi instet dies Domini 2:3 ne quis uos seducat 2:3 Let no man deceive you by any means: for that day

shall not come, except there come a falling away first, and that man of sin be revealed, the son of perdition;

ullo modo quoniam nisi uenerit discessio primum et reuelatus fuerit homo peccati filius perditionis 2:4 qui aduersatur 2:4 Who opposeth and exalteth himself above all that is

called God, or that is worshipped; so that he as God sitteth in the temple of God, shewing himself that he is God.

et extollitur supra omne quod dicitur Deus aut quod colitur ita ut in templo Dei sedeat ostendens se quia sit Deus

Biblia Sacra Vulgata

2:5 Remember ye not, that, when I was yet with you, I told 2:5 non retinetis quod cum adhuc essem apud uos haec diyou these things?

2:6 And now ye know what withholdeth that he might be cebam uobis 2:6 et nunc quid detineat scitis ut reueletur

revealed in his time.

2:7 For the mystery of iniquity doth already work: only he in suo tempore 2:7 nam mysterium iam operatur iniquitatis

who now letteth will let, until he be taken out of the way.

2:8 And then shall that Wicked be revealed, whom the tantum ut qui tenet nunc donec de medio fiat 2:8 et tunc re-

Lord shall consume with the spirit of his mouth, and shall destroy with the brightness of his coming:

uelabitur ille iniquus quem Dominus Iesus interficiet spiritu

2:9 Even him, whose coming is after the working of Satan oris sui et destruet inlustratione aduentus sui 2:9 eum cuius

with all power and signs and lying wonders,

est aduentus secundum operationem Satanae in omni uirtute

2:10 And with all deceivableness of unrighteousness in et signis et prodiiiis mendacibus 2:10 et in omni seductione

them that perish; because they received not the love of the truth, that they might be saved.

iniquitatis his qui pereunt eo quod caritatem ueritatis non re-

2:11 And for this cause God shall send them strong delu- ceperunt ut salui fierent 2:11 ideo mittit illis Deus operatio-

sion, that they should believe a lie:

2:12 That they all might be damned who believed not the nem erroris ut credant mendacio 2:12 ut iudicentur omnes

truth, but had pleasure in unrighteousness.

2:13 But we are bound to give thanks alway to God for qui non crediderunt ueritati sed consenserunt iniquitati 2:13

you, brethren beloved of the Lord, because God hath from the beginning chosen you to salvation through sanctifica-

tion of the Spirit and belief of the truth: nos autem debemus gratias agere Deo semper pro uobis frat-

res dilecti a Deo quod elegerit nos Deus primitias in salutem

2:14 Whereunto he called you by our gospel, to the ob- in sanctificatione Spiritus et fide ueritatis 2:14 ad quod et

taining of the glory of our Lord Jesus Christ.

uocauit uos per euangelium nostrum in adquisitionem glo-

2:15 Therefore, brethren, stand fast, and hold the tradi- riae Domini nostri Iesu Christi 2:15 itaque fratres state et

tions which ye have been taught, whether by word, or our epistle.

tenete traditiones quas didicistis siue per sermonem siue per

2:16 Now our Lord Jesus Christ himself, and God, even epistulam nostram 2:16 ipse autem Dominus noster Iesus

our Father, which hath loved us, and hath given us everlasting consolation and good hope through grace,

Christus et Deus et Pater noster qui dilexit nos et dedit con-

2:17 Comfort your hearts, and stablish you in every good solationem aeternam et spem bonam in gratia 2:17 exhorte-

word and work.

tur corda uestra et confirmet in omni opere et sermone bono

3:1 Finally, brethren, pray for us, that the word of the Lord 3:1 de cetero fratres orate pro nobis ut sermo Domini curmay have free course, and be glorified, even as it is with you:

rat et clarificetur sicut et apud uos 3:2 et ut liberemur ab in-

3:2 And that we may be delivered from unreasonable and wicked men: for all men have not faith. 3:3 But the Lord is faithful, who shall stablish you, and portunis et malis hominibus non enim omnium est fides 3:3 keep you from evil.

fidelis autem Dominus est qui confirmabit uos et custodiet a

3:4 And we have confidence in the Lord touching you, that malo 3:4 confidimus autem de uobis in Domino quoniam ye both do and will do the things which we command you.

3:5 And the Lord direct your hearts into the love of God, quae praecipimus et facitis et facietis 3:5 Dominus autem

and into the patient waiting for Christ.

3:6 Now we command you, brethren, in the name of our dirigat corda uestra in caritate Dei et patientia Christi 3:6

Lord Jesus Christ, that ye withdraw yourselves from every brother that walketh disorderly, and not after the tradition

which he received of us. denuntiamus autem uobis fratres in nomine Domini nostri Iesu Christi ut subtrahatis uos ab omni fratre ambulante inordinate et non secundum traditionem quam acceperunt a

3:7 For yourselves know how ye ought to follow us: for nobis 3:7 ipsi enim scitis quemadmodum oporteat imitari

we behaved not ourselves disorderly among you;

Epistola Pauli Ad Thessalonicenses II

nos quoniam non inquieti fuimus inter uos 3:8 neque gra- 3:8 Neither did we eat any man’s bread for nought; but

wrought with labour and travail night and day, that we might not be chargeable to any of you:

tis panem manducauimus ab aliquo sed in labore et fatigatione nocte et die operantes ne quem uestrum grauaremus 3:9 non quasi non habuerimus potestatem sed ut nosmet ip- 3:9 Not because we have not power, but to make ourselves

an ensample unto you to follow us.

sos formam daremus uobis ad imitandum nos 3:10 nam et 3:10 For even when we were with you, this we com-

manded you, that if any would not work, neither should he eat.

cum essemus apud uos hoc denuntiabamus uobis quoniam si quis non uult operari nec manducet 3:11 audimus enim 3:11 For we hear that there are some which walk among

you disorderly, working not at all, but are busybodies.

inter uos quosdam ambulare inquiete nihil operantes sed curiose agentes 3:12 his autem qui eiusmodi sunt denuntia- 3:12 Now them that are such we command and exhort by

our Lord Jesus Christ, that with quietness they work, and eat their own bread.

mus et obsecramus in Domino Iesu Christo ut cum silentio operantes suum panem manducent 3:13 uos autem fratres 3:13 But ye, brethren, be not weary in well doing. nolite deficere benefacientes 3:14 quod si quis non oboedit 3:14 And if any man obey not our word by this epistle,

note that man, and have no company with him, that he may be ashamed.

uerbo nostro per epistulam hunc notate et non commisceamini cum illo ut confundatur 3:15 et nolite quasi inimicum 3:15 Yet count him not as an enemy, but admonish him as

a brother.

existimare sed corripite ut fratrem 3:16 ipse autem Domi- 3:16 Now the Lord of peace himself give you peace always

by all means. The Lord be with you all.

nus pacis det uobis pacem sempiternam in omni loco Do-

minus cum omnibus uobis 3:17 salutatio mea manu Pauli 3:17 The salutation of Paul with mine own hand, which is

the token in every epistle: so I write.

quod est signum in omni epistula ita scribo 3:18 gratia Do- 3:18 The grace of our Lord Jesus Christ be with you all.

Amen.

mini nostri Iesu Christi cum omnibus uobis amen

Biblia Sacra Vulgata

MDCCVIII

Epistola Pauli Ad Timotheum I

1:1 Paulus apostolus Christi Iesu secundum imperium Dei 1:1 Paul, an apostle of Jesus Christ by the commandment

of God our Saviour, and Lord Jesus Christ, which is our hope;

saluatoris nostri et Christi Iesu spei nostrae 1:2 Timotheo

1:2 Unto Timothy, my own son in the faith: Grace, mercy, and peace, from God our Father and Jesus Christ our Lord.

dilecto filio in fide gratia misericordia pax a Deo Patre et Christo Iesu Domino nostro 1:3 sicut rogaui te ut remaneres 1:3 As I besought thee to abide still at Ephesus, when I

went into Macedonia, that thou mightest charge some that they teach no other doctrine,

Ephesi cum irem in Macedoniam ut denuntiares quibusdam ne aliter docerent 1:4 neque intenderent fabulis et genea- 1:4 Neither give heed to fables and endless genealo-

gies, which minister questions, rather than godly edifying which is in faith: so do.

logiis interminatis quae quaestiones praestant magis quam aedificationem Dei quae est in fide 1:5 finis autem prae- 1:5 Now the end of the commandment is charity out of

a pure heart, and of a good conscience, and of faith unfeigned:

cepti est caritas de corde puro et conscientia bona et fide non ficta 1:6 a quibus quidam aberrantes conuersi sunt in uani- 1:6 From which some having swerved have turned aside

unto vain jangling;

loquium 1:7 uolentes esse legis doctores non intellegentes 1:7 Desiring to be teachers of the law; understanding nei-

ther what they say, nor whereof they affirm.

neque quae loquuntur neque de quibus adfirmant 1:8 sci- 1:8 But we know that the law is good, if a man use it law-

fully;

mus autem quia bona est lex si quis ea legitime utatur 1:9 1:9 Knowing this, that the law is not made for a righteous

man, but for the lawless and disobedient, for the ungodly and for sinners, for unholy and profane, for murderers of

sciens hoc quia iusto lex non est posita sed iniustis et non fathers and murderers of mothers, for manslayers, subditis impiis et peccatoribus sceleratis et contaminatis patricidis et matricidis homicidis 1:10 fornicariis masculorum 1:10 For whoremongers, for them that defile themselves

with mankind, for menstealers, for liars, for perjured persons, and if there be any other thing that is contrary to

concubitoribus plagiariis mendacibus periuris et si quid aliud sound doctrine; sanae doctrinae aduersatur 1:11 quae est secundum euan- 1:11 According to the glorious gospel of the blessed God,

which was committed to my trust.

gelium gloriae beati Dei quod creditum est mihi 1:12 gra- 1:12 And I thank Christ Jesus our Lord, who hath enabled

me, for that he counted me faithful, putting me into the ministry;

tias ago ei qui me confortauit Christo Iesu Domino nostro quia fidelem me existimauit ponens in ministerio 1:13 qui 1:13 Who was before a blasphemer, and a persecutor, and

injurious: but I obtained mercy, because I did it ignorantly in unbelief.

prius fui blasphemus et persecutor et contumeliosus sed misericordiam consecutus sum quia ignorans feci in incredulitate 1:14 superabundauit autem gratia Domini nostri cum 1:14 And the grace of our Lord was exceeding abundant

with faith and love which is in Christ Jesus.

fide et dilectione quae est in Christo Iesu 1:15 fidelis sermo 1:15 This is a faithful saying, and worthy of all accepta-

tion, that Christ Jesus came into the world to save sinners; of whom I am chief.

et omni acceptione dignus quia Christus Iesus uenit in mundum peccatores saluos facere quorum primus ego sum 1:16 1:16 Howbeit for this cause I obtained mercy, that in me

first Jesus Christ might shew forth all longsuffering, for a pattern to them which should hereafter believe on him to

sed ideo misericordiam consecutus sum ut in me primo os- life everlasting. tenderet Christus Iesus omnem patientiam ad deformatio-

nem eorum qui credituri sunt illi in uitam aeternam 1:17 1:17 Now unto the King eternal, immortal, invisible, the only wise God, be honour and glory for ever and ever. Amen.

regi autem saeculorum inmortali inuisibili soli Deo honor

et gloria in saecula saeculorum amen 1:18 hoc praecep- 1:18 This charge I commit unto thee, son Timothy, according to the prophecies which went before on thee, that thou by them mightest war a good warfare;

tum commendo tibi fili Timothee secundum praecedentes 1:19 Holding faith, and a good conscience; which some in te prophetias ut milites in illis bonam militiam 1:19 ha-

having put away concerning faith have made shipwreck:

bens fidem et bonam conscientiam quam quidam repellentes

1:20 Of whom is Hymenaeus and Alexander; whom I circa fidem naufragauerunt 1:20 ex quibus est Hymeneus et

have delivered unto Satan, that they may learn not to blaspheme.

Alexander quos tradidi Satanae ut discant non blasphemare

2:1 I exhort therefore, that, first of all, supplications, 2:1 obsecro igitur primo omnium fieri obsecrationes oraprayers, intercessions, and giving of thanks, be made for all men;

tiones postulationes gratiarum actiones pro omnibus homi-

2:2 For kings, and for all that are in authority; that we may nibus 2:2 pro regibus et omnibus qui in sublimitate sunt lead a quiet and peaceable life in all godliness and honesty.

ut quietam et tranquillam uitam agamus in omni pietate et

2:3 For this is good and acceptable in the sight of God our castitate 2:3 hoc enim bonum est et acceptum coram sa-Saviour;

2:4 Who will have all men to be saved, and to come unto lutari nostro Deo 2:4 qui omnes homines uult saluos fieri

the knowledge of the truth.

2:5 For there is one God, and one mediator between God et ad agnitionem ueritatis uenire 2:5 unus enim Deus unus

and men, the man Christ Jesus; 2:6 Who gave himself a ransom for all, to be testified in et mediator Dei et hominum homo Christus Iesus 2:6 qui due time.

dedit redeeptionem semet ipsum pro omnibus testimonium

2:7 Whereunto I am ordained a preacher, and an apostle, temporibus suis 2:7 in quo positus sum ego praedicator et

(I speak the truth in Christ, and lie not;) a teacher of the Gentiles in faith and verity.

apostolus ueritatem dico non mentior doctor gentium in fide

2:8 I will therefore that men pray every where, lifting up et ueritate 2:8 uolo ergo uiros orare in omni loco leuantes

holy hands, without wrath and doubting.

2:9 In like manner also, that women adorn themselves in puras manus sine ira et disceptatione 2:9 similiter et mu-

modest apparel, with shamefacedness and sobriety; not with broided hair, or gold, or pearls, or costly array;

lieres in habitu ornato cum uerecundia et sobrietate ornantes se non in tortis crinibus aut auro aut margaritis uel ueste

2:10 But (which becometh women professing godliness) pretiosa 2:10 sed quod decet mulieres promittentes pieta-

with good works.

2:11 Let the woman learn in silence with all subjection. tem per opera bona 2:11 mulier in silentio discat cum omni 2:12 But I suffer not a woman to teach, nor to usurp au- subiectione 2:12 docere autem mulieri non permitto neque

thority over the man, but to be in silence.

2:13 For Adam was first formed, then Eve. dominari in uirum sed esse in silentio 2:13 Adam enim pri-2:14 And Adam was not deceived, but the woman being mus formatus est deinde Eua 2:14 et Adam non est seductus

deceived was in the transgression.

2:15 Notwithstanding she shall be saved in childbearing, mulier autem seducta in praeuaricatione fuit 2:15 saluabi-

if they continue in faith and charity and holiness with sobriety.

tur autem per filiorum generationem si permanserint in fide et dilectione et sanctificatione cum sobrietate

3:1 This is a true saying, If a man desire the office of a 3:1 fidelis sermo si quis episcopatum desiderat bonum

bishop, he desireth a good work.

3:2 A bishop then must be blameless, the husband of one opus desiderat 3:2 oportet ergo episcopum inreprehensibi-

wife, vigilant, sober, of good behaviour, given to hospitality, apt to teach;

lem esse unius uxoris uirum sobrium prudentem ornatum

3:3 Not given to wine, no striker, not greedy of filthy lucre; hospitalem doctorem 3:3 non uinolentum non percussorem but patient, not a brawler, not covetous; 3:4 One that ruleth well his own house, having his children sed modestum non litigiosum non cupidum 3:4 suae domui in subjection with all gravity;

bene praepositum filios habentem subditos cum omni casti-

Epistola Pauli Ad Timotheum I

tate 3:5 si quis autem domui suae praeesse nescit quomodo 3:5 (For if a man know not how to rule his own house, how

shall he take care of the church of God?)

ecclesiae Dei diligentiam habebit 3:6 non neophytum ne in 3:6 Not a novice, lest being lifted up with pride he fall into

the condemnation of the devil.

superbia elatus in iudicium incidat diaboli 3:7 oportet au- 3:7 Moreover he must have a good report of them which

are without; lest he fall into reproach and the snare of the devil.

tem illum et testimonium habere bonum ab his qui foris sunt ut non in obprobrium incidat et laqueum diaboli 3:8 diaco- 3:8 Likewise must the deacons be grave, not double-

tongued, not given to much wine, not greedy of filthy lucre;

nos similiter pudicos non bilingues non multo uino deditos

non turpe lucrum sectantes 3:9 habentes mysterium fidei 3:9 Holding the mystery of the faith in a pure conscience.

in conscientia pura 3:10 et hii autem probentur primum et 3:10 And let these also first be proved; then let them use

the office of a deacon, being found blameless.

sic ministrent nullum crimen habentes 3:11 mulieres simili- 3:11 Even so must their wives be grave, not slanderers,

sober, faithful in all things.

ter pudicas non detrahentes sobrias fideles in omnibus 3:12 3:12 Let the deacons be the husbands of one wife, ruling

their children and their own houses well.

diacones sint unius uxoris uiri qui filiis suis bene praesunt et suis domibus 3:13 qui enim bene ministrauerint gradum 3:13 For they that have used the office of a deacon well

purchase to themselves a good degree, and great boldness in the faith which is in Christ Jesus.

sibi bonum adquirent et multam fiduciam in fide quae est in Christo Iesu 3:14 haec tibi scribo sperans uenire ad te cito 3:14 These things write I unto thee, hoping to come unto

thee shortly:

3:15 si autem tardauero ut scias quomodo oporteat te in domo 3:15 But if I tarry long, that thou mayest know how thou

oughtest to behave thyself in the house of God, which is the church of the living God, the pillar and ground of the

Dei conuersari quae est ecclesia Dei uiui columna et firma- truth. mentum ueritatis 3:16 et manifeste magnum est pietatis sa- 3:16 And without controversy great is the mystery of god-

liness: God was manifest in the flesh, justified in the Spirit, seen of angels, preached unto the Gentiles, believed on in

cramentum quod manifestatum est in carne iustificatum est the world, received up into glory. in spiritu apparuit angelis praedicatum est gentibus creditum est in mundo adsumptum est in gloria

4:1 Spiritus autem manifeste dicit quia in nouissimis tem- 4:1 Now the Spirit speaketh expressly, that in the latter

times some shall depart from the faith, giving heed to seducing spirits, and doctrines of devils;

poribus discedent quidam a fide adtendentes spiritibus erroris et doctrinis daemoniorum 4:2 in hypocrisi loquen- 4:2 Speaking lies in hypocrisy; having their conscience

seared with a hot iron;

tium mendacium et cauteriatam habentium suam conscientiam 4:3 prohibentium nubere abstinere a cibis quos De- 4:3 Forbidding to marry, and commanding to abstain from

meats, which God hath created to be received with thanksgiving of them which believe and know the truth.

us creauit ad percipiendum cum gratiarum actione fidelibus et his qui cognouerunt ueritatem 4:4 quia omnis creatura 4:4 For every creature of God is good, and nothing to be

refused, if it be received with thanksgiving:

Dei bona et nihil reiciendum quod cum gratiarum actione percipitur 4:5 sanctificatur enim per uerbum Dei et oratio- 4:5 For it is sanctified by the word of God and prayer. nem 4:6 haec proponens fratribus bonus eris minister Ch- 4:6 If thou put the brethren in remembrance of these

things, thou shalt be a good minister of Jesus Christ, nourished up in the words of faith and of good doctrine, where-

risti Iesu enutritus uerbis fidei et bonae doctrinae quam adse- unto thou hast attained. cutus es 4:7 ineptas autem et aniles fabulas deuita exerce te 4:7 But refuse profane and old wives’ fables, and exercise

thyself rather unto godliness.

ipsum ad pietatem 4:8 nam corporalis exercitatio ad modi- 4:8 For bodily exercise profiteth little: but godliness is

profitable unto all things, having promise of the life that now is, and of that which is to come.

Biblia Sacra Vulgata

cum utilis est pietas autem ad omnia utilis est promissionem

4:9 This is a faithful saying and worthy of all acceptation. habens uitae quae nunc est et futurae 4:9 fidelis sermo et 4:10 For therefore we both labour and suffer reproach, be- omni acceptione dignus 4:10 in hoc enim laboramus et macause we trust in the living God, who is the Saviour of all men, specially of those that believe.

ledicimur quia sperauimus in Deum uiuum qui est saluator

4:11 These things command and teach. omnium hominum maxime fidelium 4:11 praecipe haec et

4:12 Let no man despise thy youth; but be thou an example doce 4:12 nemo adulescentiam tuam contemnat sed exemof the believers, in word, in conversation, in charity, in spirit, in faith, in purity.

plum esto fidelium in uerbo in conuersatione in caritate in

4:13 Till I come, give attendance to reading, to exhorta- fide in castitate 4:13 dum uenio adtende lectioni exhorta-

tion, to doctrine.

4:14 Neglect not the gift that is in thee, which was given tioni doctrinae 4:14 noli neglegere gratiam quae in te est

thee by prophecy, with the laying on of the hands of the presbytery.

quae data est tibi per prophetiam cum inpositione manuum

4:15 Meditate upon these things; give thyself wholly to presbyterii 4:15 haec meditare in his esto ut profectus tuus

them; that thy profiting may appear to all.

4:16 Take heed unto thyself, and unto the doctrine; con- manifestus sit omnibus 4:16 adtende tibi et doctrinae insta

tinue in them: for in doing this thou shalt both save thyself, and them that hear thee.

in illis hoc enim faciens et te ipsum saluum facies et qui te audiunt

5:1 Rebuke not an elder, but intreat him as a father; and 5:1 seniorem ne increpaueris sed obsecra ut patrem iuthe younger men as brethren;

5:2 The elder women as mothers; the younger as sisters, uenes ut fratres 5:2 anus ut matres iuuenculas ut sorores

with all purity.

5:3 Honour widows that are widows indeed. in omni castitate 5:3 uiduas honora quae uere uiduae sunt 5:4 But if any widow have children or nephews, let them 5:4 si qua autem uidua filios aut nepotes habet discant pri-

learn first to shew piety at home, and to requite their parents: for that is good and acceptable before God.

mum domum suam regere et mutuam uicem reddere parenti-

5:5 Now she that is a widow indeed, and desolate, trusteth bus hoc enim acceptum est coram Deo 5:5 quae autem uere in God, and continueth in supplications and prayers night and day.

uidua est et desolata sperauit in Deum et instat obsecratio-

5:6 But she that liveth in pleasure is dead while she liveth. nibus et orationibus nocte ac die 5:6 nam quae in deliciis

5:7 And these things give in charge, that they may be est uiuens mortua est 5:7 et hoc praecipe ut inreprehensi-

blameless.

5:8 But if any provide not for his own, and specially for biles sint 5:8 si quis autem suorum et maxime domestico-

those of his own house, he hath denied the faith, and is worse than an infidel.

rum curam non habet fidem negauit et est infideli deterior

5:9 Let not a widow be taken into the number under three- 5:9 uidua eligatur non minus sexaginta annorum quae fuerit

score years old, having been the wife of one man.

5:10 Well reported of for good works; if she have brought unius uiri uxor 5:10 in operibus bonis testimonium habens

up children, if she have lodged strangers, if she have washed the saints’ feet, if she have relieved the afflicted,

if she have diligently followed every good work. si filios educauit si hospitio recepit si sanctorum pedes lauit

si tribulationem patientibus subministrauit si omne opus bo-

5:11 But the younger widows refuse: for when they have num subsecuta est 5:11 adulescentiores autem uiduas de-

begun to wax wanton against Christ, they will marry;

uita cum enim luxuriatae fuerint in Christo nubere uolunt

5:12 Having damnation, because they have cast off their 5:12 habentes damnationem quia primam fidem irritam fece-

first faith.

5:13 And withal they learn to be idle, wandering about runt 5:13 simul autem et otiosae discunt circumire domos

from house to house; and not only idle, but tattlers also and busybodies, speaking things which they ought not.

Epistola Pauli Ad Timotheum I

non solum otiosae sed et uerbosae et curiosae loquentes quae non oportet 5:14 uolo ergo iuueniores nubere filios procre- 5:14 I will therefore that the younger women marry, bear

children, guide the house, give none occasion to the adversary to speak reproachfully.

are matres familias esse nullam occasionem dare aduersario maledicti gratia 5:15 iam enim quaedam conuersae sunt re- 5:15 For some are already turned aside after Satan. tro Satanan 5:16 si qua fidelis habet uiduas subministret il- 5:16 If any man or woman that believeth have widows, let

them relieve them, and let not the church be charged; that it may relieve them that are widows indeed.

lis et non grauetur ecclesia ut his quae uere uiduae sunt sufficiat 5:17 qui bene praesunt presbyteri duplici honore digni 5:17 Let the elders that rule well be counted worthy of

double honour, especially they who labour in the word and doctrine.

habeantur maxime qui laborant in uerbo et doctrina 5:18 di-

5:18 For the scripture saith, Thou shalt not muzzle the ox that treadeth out the corn. And, The labourer is worthy of his reward.

cit enim scriptura non infrenabis os boui trituranti et dignus

operarius mercede sua 5:19 aduersus presbyterum accusa- 5:19 Against an elder receive not an accusation, but before

two or three witnesses.

tionem noli recipere nisi sub duobus et tribus testibus 5:20 5:20 Them that sin rebuke before all, that others also may

fear.

peccantes coram omnibus argue ut et ceteri timorem habeant 5:21 testor coram Deo et Christo Iesu et electis angelis ut 5:21 I charge thee before God, and the Lord Jesus Christ,

and the elect angels, that thou observe these things without preferring one before another, doing nothing by partiality.

haec custodias sine praeiudicio nihil faciens in aliam partem declinando 5:22 manus cito nemini inposueris neque com- 5:22 Lay hands suddenly on no man, neither be partaker

of other men’s sins: keep thyself pure.

municaueris peccatis alienis te ipsum castum custodi 5:23 5:23 Drink no longer water, but use a little wine for thy

stomach’s sake and thine often infirmities.

noli adhuc aquam bibere sed uino modico utere propter stomachum tuum et frequentes tuas infirmitates 5:24 quorun- 5:24 Some men’s sins are open beforehand, going before

to judgment; and some men they follow after.

dam hominum peccata manifesta sunt praecedentia ad iudicium quosdam autem et subsequuntur 5:25 similiter et facta 5:25 Likewise also the good works of some are manifest

beforehand; and they that are otherwise cannot be hid.

bona manifesta sunt et quae aliter se habent abscondi non possunt

6:1 quicumque sunt sub iugo serui dominos suos omni ho- 6:1 Let as many servants as are under the yoke count their own masters worthy of all honour, that the name of God and his doctrine be not blasphemed.

nore dignos arbitrentur ne nomen Domini et doctrina blas-

phemetur 6:2 qui autem fideles habent dominos non con- 6:2 And they that have believing masters, let them not despise them, because they are brethren; but rather do them service, because they are faithful and beloved, partakers of

temnant quia fratres sunt sed magis seruiant quia fideles sunt the benefit. These things teach and exhort. et dilecti qui beneficii participes sunt haec doce et exhortare 6:3 si quis aliter docet et non adquiescit sanis sermoni- 6:3 If any man teach otherwise, and consent not to whole-

some words, even the words of our Lord Jesus Christ, and to the doctrine which is according to godliness;

bus Domini nostri Iesu Christi et ei quae secundum pietatem est doctrinae 6:4 superbus nihil sciens sed languens circa 6:4 He is proud, knowing nothing, but doting about ques-

tions and strifes of words, whereof cometh envy, strife, railings, evil surmisings,

quaestiones et pugnas uerborum ex quibus oriuntur inuidiae contentiones blasphemiae suspiciones malae 6:5 conflicta- 6:5 Perverse disputings of men of corrupt minds, and des-

titute of the truth, supposing that gain is godliness: from such withdraw thyself.

tiones hominum mente corruptorum et qui ueritate priuati

Biblia Sacra Vulgata

6:6 But godliness with contentment is great gain. sunt existimantium quaestum esse pietatem 6:6 est autem

6:7 For we brought nothing into this world, and it is certain quaestus magnus pietas cum sufficientia 6:7 nihil enim inwe can carry nothing out.

tulimus in mundum haut dubium quia nec auferre quid pos-

6:8 And having food and raiment let us be therewith con- sumus 6:8 habentes autem alimenta et quibus tegamur his

tent.

6:9 But they that will be rich fall into temptation and contenti sumus 6:9 nam qui uolunt diuites fieri incidunt

a snare, and into many foolish and hurtful lusts, which drown men in destruction and perdition.

in temptationem et laqueum et desideria multa inutilia et nociua quae mergunt homines in interitum et perditionem

6:10 For the love of money is the root of all evil: which 6:10 radix enim omnium malorum est cupiditas quam qui-

while some coveted after, they have erred from the faith, and pierced themselves through with many sorrows.

dam appetentes errauerunt a fide et inseruerunt se doloribus

6:11 But thou, O man of God, flee these things; and follow multis 6:11 tu autem o homo Dei haec fuge sectare uero iu-

after righteousness, godliness, faith, love, patience, meekness.

stitiam pietatem fidem caritatem patientiam mansuetudinem

6:12 Fight the good fight of faith, lay hold on eternal life, 6:12 certa bonum certamen fidei adprehende uitam aeternam whereunto thou art also called, and hast professed a good profession before many witnesses.

in qua uocatus es et confessus bonam confessionem coram

6:13 I give thee charge in the sight of God, who quickeneth multis testibus 6:13 praecipio tibi coram Deo qui uiuificat

all things, and before Christ Jesus, who before Pontius Pilate witnessed a good confession;

omnia et Christo Iesu qui testimonium reddidit sub Pontio

6:14 That thou keep this commandment without spot, un- Pilato bonam confessionem 6:14 ut serues mandatum sine

rebukable, until the appearing of our Lord Jesus Christ:

macula inreprehensibile usque in aduentum Domini nostri

6:15 Which in his times he shall shew, who is the blessed Iesu Christi 6:15 quem suis temporibus ostendet beatus et

and only Potentate, the King of kings, and Lord of lords;

6:16 Who only hath immortality, dwelling in the light solus potens rex regum et Dominus dominantium 6:16 qui

which no man can approach unto; whom no man hath seen, nor can see: to whom be honour and power ever-

lasting. Amen. solus habet inmortalitatem lucem habitans inaccessibilem

quem uidit nullus hominum sed nec uidere potest cui honor

6:17 Charge them that are rich in this world, that they be et imperium sempiternum amen 6:17 diuitibus huius saenot highminded, nor trust in uncertain riches, but in the living God, who giveth us richly all things to enjoy;

culi praecipe non sublime sapere neque sperare in incerto diuitiarum sed in Deo qui praestat nobis omnia abunde ad

6:18 That they do good, that they be rich in good works, fruendum 6:18 bene agere diuites fieri in operibus bonis

ready to distribute, willing to communicate;

6:19 Laying up in store for themselves a good foundation facile tribuere communicare 6:19 thesaurizare sibi funda-

against the time to come, that they may lay hold on eternal life.

mentum bonum in futurum ut adprehendant ueram uitam

6:20 O Timothy, keep that which is committed to thy trust, 6:20 o Timothee depositum custodi deuitans profanas uocum

avoiding profane and vain babblings, and oppositions of science falsely so called:

nouitates et oppositiones falsi nominis scientiae 6:21 quam

6:21 Which some professing have erred concerning the faith. Grace be with thee. Amen.

quidam promittentes circa fidem exciderunt gratia tecum

Epistola Pauli Ad Timotheum II

1:1 Paulus apostolus Christi Iesu per uoluntatem Dei se- 1:1 Paul, an apostle of Jesus Christ by the will of God,

according to the promise of life which is in Christ Jesus,

cundum promissionem uitae quae est in Christo Iesu 1:2 1:2 To Timothy, my dearly beloved son: Grace, mercy, and

peace, from God the Father and Christ Jesus our Lord.

Timotheo carissimo filio gratia misericordia pax a Deo Patre et Christo Iesu Domino nostro 1:3 gratias ago Deo cui 1:3 I thank God, whom I serve from my forefathers with

pure conscience, that without ceasing I have remembrance of thee in my prayers night and day;

seruio a progenitoribus in conscientia pura quam sine intermissione habeam tui memoriam in orationibus meis nocte ac

die 1:4 desiderans te uidere memor lacrimarum tuarum ut 1:4 Greatly desiring to see thee, being mindful of thy tears,

that I may be filled with joy;

gaudio implear 1:5 recordationem accipiens eius fidei quae 1:5 When I call to remembrance the unfeigned faith that is

in thee, which dwelt first in thy grandmother Lois, and thy mother Eunice; and I am persuaded that in thee also.

est in te non ficta quae et habitauit primum in auia tua Lo-

ide et matre tua Eunice certus sum autem quod et in te 1:6 1:6 Wherefore I put thee in remembrance that thou stir up the gift of God, which is in thee by the putting on of my hands.

propter quam causam admoneo te ut resuscites gratiam Dei quae est in te per inpositionem manuum mearum 1:7 non 1:7 For God hath not given us the spirit of fear; but of

power, and of love, and of a sound mind.

enim dedit nobis Deus spiritum timoris sed uirtutis et dilectionis et sobrietatis 1:8 noli itaque erubescere testimo- 1:8 Be not thou therefore ashamed of the testimony of our

Lord, nor of me his prisoner: but be thou partaker of the afflictions of the gospel according to the power of God;

nium Domini nostri neque me uinctum eius sed conlabora euangelio secundum uirtutem Dei 1:9 qui nos liberauit et 1:9 Who hath saved us, and called us with an holy calling,

not according to our works, but according to his own purpose and grace, which was given us in Christ Jesus before

uocauit uocatione sancta non secundum opera nostra sed se- the world began, cundum propositum suum et gratiam quae data est nobis in Christo Iesu ante tempora saecularia 1:10 manifestata est 1:10 But is now made manifest by the appearing of our

Saviour Jesus Christ, who hath abolished death, and hath brought life and immortality to light through the gospel:

autem nunc per inluminationem saluatoris nostri Iesu Christi qui destruxit quidem mortem inluminauit autem uitam et in-

corruptionem per euangelium 1:11 in quo positus sum ego 1:11 Whereunto I am appointed a preacher, and an apostle,

and a teacher of the Gentiles.

praedicator et apostolus et magister gentium 1:12 ob quam 1:12 For the which cause I also suffer these things: nev-

ertheless I am not ashamed: for I know whom I have believed, and am persuaded that he is able to keep that which

causam etiam haec patior sed non confundor scio enim cui I have committed unto him against that day. credidi et certus sum quia potens est depositum meum ser-

uare in illum diem 1:13 formam habe sanorum uerborum 1:13 Hold fast the form of sound words, which thou hast

heard of me, in faith and love which is in Christ Jesus.

quae a me audisti in fide et dilectione in Christo Iesu 1:14 1:14 That good thing which was committed unto thee keep

by the Holy Ghost which dwelleth in us.

bonum depositum custodi per Spiritum Sanctum qui habi-

tat in nobis 1:15 scis hoc quod auersi sunt a me omnes qui 1:15 This thou knowest, that all they which are in Asia be turned away from me; of whom are Phygellus and Hermogenes.

in Asia sunt ex quibus est Phygelus et Hermogenes 1:16

1:16 The Lord give mercy unto the house of Onesiphorus; for he oft refreshed me, and was not ashamed of my chain:

det misericordiam Dominus Onesifori domui quia saepe me refrigerauit et catenam meam non erubuit 1:17 sed cum 1:17 But, when he was in Rome, he sought me out very

diligently, and found me.

Romam uenisset sollicite me quaesiuit et inuenit 1:18 det 1:18 The Lord grant unto him that he may find mercy of

the Lord in that day: and in how many things he ministered unto me at Ephesus, thou knowest very well.

illi Dominus inuenire misericordiam a Domino in illa die et quanta Ephesi ministrauit melius tu nosti

2:1 Thou therefore, my son, be strong in the grace that is 2:1 tu ergo fili mi confortare in gratia quae est in Christo

in Christ Jesus.

2:2 And the things that thou hast heard of me among many Iesu 2:2 et quae audisti a me per multos testes haec com-

witnesses, the same commit thou to faithful men, who shall be able to teach others also.

menda fidelibus hominibus qui idonei erunt et alios docere

2:3 Thou therefore endure hardness, as a good soldier of 2:3 labora sicut bonus miles Christi Iesu 2:4 nemo militans

Jesus Christ.

2:4 No man that warreth entangleth himself with the affairs of this life; that he may please him who hath chosen him to be a soldier.

inplicat se negotiis saecularibus ut ei placeat cui se probauit

2:5 And if a man also strive for masteries, yet is he not 2:5 nam et qui certat in agone non coronatur nisi legitime cer-

crowned, except he strive lawfully.

2:6 The husbandman that laboureth must be first partaker tauerit 2:6 laborantem agricolam oportet primum de fruc-

of the fruits.

2:7 Consider what I say; and the Lord give thee under- tibus accipere 2:7 intellege quae dico dabit enim tibi Do-

standing in all things.

2:8 Remember that Jesus Christ of the seed of David was minus in omnibus intellectum 2:8 memor esto Iesum Chraised from the dead according to my gospel:

ristum resurrexisse a mortuis ex semine Dauid secundum

2:9 Wherein I suffer trouble, as an evil doer, even unto euangelium meum 2:9 in quo laboro usque ad uincula quasi

bonds; but the word of God is not bound.

2:10 Therefore I endure all things for the elect’s sakes, that male operans sed uerbum Dei non est alligatum 2:10 ideo they may also obtain the salvation which is in Christ Jesus with eternal glory.

omnia sustineo propter electos ut et ipsi salutem consequan-

2:11 It is a faithful saying: For if we be dead with him, we tur quae est in Christo Iesu cum gloria caelesti 2:11 fidelis

shall also live with him:

2:12 If we suffer, we shall also reign with him: if we deny sermo nam si conmortui sumus et conuiuemus 2:12 si suhim, he also will deny us:

stinemus et conregnabimus si negabimus et ille negabit nos

2:13 If we believe not, yet he abideth faithful: he cannot 2:13 si non credimus ille fidelis manet negare se ipsum non deny himself.

2:14 Of these things put them in remembrance, charging potest 2:14 haec commone testificans coram Domino noli them before the Lord that they strive not about words to no profit, but to the subverting of the hearers.

uerbis contendere in nihil utile ad subuersionem audientium

2:15 Study to shew thyself approved unto God, a workman 2:15 sollicite cura te ipsum probabilem exhibere Deo opera-

that needeth not to be ashamed, rightly dividing the word of truth.

rium inconfusibilem recte tractantem uerbum ueritatis 2:16

2:16 But shun profane and vain babblings: for they will increase unto more ungodliness.

profana autem inaniloquia deuita multum enim proficient ad

2:17 And their word will eat as doth a canker: of whom is impietatem 2:17 et sermo eorum ut cancer serpit ex quibus

Hymenaeus and Philetus;

2:18 Who concerning the truth have erred, saying that the est Hymeneus et Philetus 2:18 qui a ueritate exciderunt di-

resurrection is past already; and overthrow the faith of some.

centes resurrectionem iam factam et subuertunt quorundam

2:19 Nevertheless the foundation of God standeth sure, fidem 2:19 sed firmum fundamentum Dei stetit habens sighaving this seal, The Lord knoweth them that are his. And, Let every one that nameth the name of Christ depart from

iniquity. naculum hoc cognouit Dominus qui sunt eius et discedat ab

2:20 But in a great house there are not only vessels of gold iniquitate omnis qui nominat nomen Domini 2:20 in magna and of silver, but also of wood and of earth; and some to honour, and some to dishonour.

autem domo non solum sunt uasa aurea et argentea sed et lignea et fictilia et quaedam quidem in honorem quaedam

2:21 If a man therefore purge himself from these, he shall autem in contumeliam 2:21 si quis ergo emundauerit se ab be a vessel unto honour, sanctified, and meet for the master’s use, and prepared unto every good work.

Epistola Pauli Ad Timotheum II

istis erit uas in honorem sanctificatum et utile Domino ad

omne opus bonum paratum 2:22 iuuenilia autem desideria 2:22 Flee also youthful lusts: but follow righteousness, faith, charity, peace, with them that call on the Lord out of a pure heart.

fuge sectare uero iustitiam fidem caritatem pacem cum his

qui inuocant Dominum de corde puro 2:23 stultas autem et 2:23 But foolish and unlearned questions avoid, knowing

that they do gender strifes.

sine disciplina quaestiones deuita sciens quia generant lites 2:24 seruum autem Domini non oportet litigare sed mansue- 2:24 And the servant of the Lord must not strive; but be

gentle unto all men, apt to teach, patient,

tum esse ad omnes docibilem patientem 2:25 cum modestia 2:25 In meekness instructing those that oppose them-

selves; if God peradventure will give them repentance to the acknowledging of the truth;

corripientem eos qui resistunt nequando det illis Deus paenitentiam ad cognoscendam ueritatem 2:26 et resipiscant a 2:26 And that they may recover themselves out of the

snare of the devil, who are taken captive by him at his will.

diaboli laqueis a quo capti tenentur ad ipsius uoluntatem

3:1 hoc autem scito quod in nouissimis diebus instabunt 3:1 This know also, that in the last days perilous times

shall come.

tempora periculosa 3:2 et erunt homines se ipsos amantes 3:2 For men shall be lovers of their own selves, covetous,

boasters, proud, blasphemers, disobedient to parents, unthankful, unholy,

cupidi elati superbi blasphemi parentibus inoboedientes ingrati scelesti 3:3 sine affectione sine pace criminatores in- 3:3 Without natural affection, trucebreakers, false ac-

cusers, incontinent, fierce, despisers of those that are good,

continentes inmites sine benignitate 3:4 prodiiores proterui 3:4 Traitors, heady, highminded, lovers of pleasures more

than lovers of God;

tumidi uoluptatium amatores magis quam Dei 3:5 haben- 3:5 Having a form of godliness, but denying the power

thereof: from such turn away.

tes speciem quidem pietatis uirtutem autem eius abnegantes

et hos deuita 3:6 ex his enim sunt qui penetrant domos et 3:6 For of this sort are they which creep into houses, and lead captive silly women laden with sins, led away with divers lusts,

captiuas ducunt mulierculas oneratas peccatis quae ducuntur uariis desideriis 3:7 semper discentes et numquam ad 3:7 Ever learning, and never able to come to the knowl-

edge of the truth.

scientiam ueritatis peruenientes 3:8 quemadmodum autem 3:8 Now as Jannes and Jambres withstood Moses, so do

these also resist the truth: men of corrupt minds, reprobate concerning the faith.

Iannes et Mambres restiterunt Mosi ita et hii resistunt ueritati

homines corrupti mente reprobi circa fidem 3:9 sed ultra 3:9 But they shall proceed no further: for their folly shall

be manifest unto all men, as their’s also was.

non proficient insipientia enim eorum manifesta erit omnibus sicut et illorum fuit 3:10 tu autem adsecutus es meam 3:10 But thou hast fully known my doctrine, manner of

life, purpose, faith, longsuffering, charity, patience,

doctrinam institutionem propositum fidem longanimitatem dilectionem patientiam 3:11 persecutiones passiones qualia 3:11 Persecutions, afflictions, which came unto me at An-

tioch, at Iconium, at Lystra; what persecutions I endured: but out of them all the Lord delivered me.

mihi facta sunt Antiochiae Iconii Lystris quales persecutio-

nes sustinui et ex omnibus me eripuit Dominus 3:12 et om- 3:12 Yea, and all that will live godly in Christ Jesus shall

suffer persecution.

nes qui uolunt pie uiuere in Christo Iesu persecutionem patientur 3:13 mali autem homines et seductores proficient in 3:13 But evil men and seducers shall wax worse and

worse, deceiving, and being deceived.

peius errantes et in errorem mittentes 3:14 tu uero permane 3:14 But continue thou in the things which thou hast

learned and hast been assured of, knowing of whom thou hast learned them;

in his quae didicisti et credita sunt tibi sciens a quo didiceris

Biblia Sacra Vulgata

3:15 And that from a child thou hast known the holy scrip- 3:15 et quia ab infantia sacras litteras nosti quae te possint

tures, which are able to make thee wise unto salvation through faith which is in Christ Jesus.

instruere ad salutem per fidem quae est in Christo Iesu 3:16

3:16 All scripture is given by inspiration of God, and is profitable for doctrine, for reproof, for correction, for instruction in righteousness:

omnis scriptura diuinitus inspirata et utilis ad docendum ad

3:17 That the man of God may be perfect, throughly fur- arguendum ad corrigendum ad erudiendum in iustitia 3:17

nished unto all good works.

ut perfectus sit homo Dei ad omne opus bonum instructus

4:1 I charge thee therefore before God, and the Lord Je- 4:1 testificor coram Deo et Christo Iesu qui iudicaturus

sus Christ, who shall judge the quick and the dead at his appearing and his kingdom;

est uiuos ac mortuos et aduentum ipsius et regnum eius 4:2

4:2 Preach the word; be instant in season, out of season; reprove, rebuke, exhort with all longsuffering and doctrine.

praedica uerbum insta oportune inportune argue obsecra in-

4:3 For the time will come when they will not endure crepa in omni patientia et doctrina 4:3 erit enim tempus

sound doctrine; but after their own lusts shall they heap to themselves teachers, having itching ears;

cum sanam doctrinam non sustinebunt sed ad sua deside-

4:4 And they shall turn away their ears from the truth, and ria coaceruabunt sibi magistros prurientes auribus 4:4 et a

shall be turned unto fables.

ueritate quidem auditum auertent ad fabulas autem conuer-

4:5 But watch thou in all things, endure afflictions, do the tentur 4:5 tu uero uigila in omnibus labora opus fac euanwork of an evangelist, make full proof of thy ministry.

4:6 For I am now ready to be offered, and the time of my gelistae ministerium tuum imple 4:6 ego enim iam deli-

departure is at hand.

4:7 I have fought a good fight, I have finished my course, bor et tempus meae resolutionis instat 4:7 bonum certamen

I have kept the faith:

4:8 Henceforth there is laid up for me a crown of righ- certaui cursum consummaui fidem seruaui 4:8 in reliquo

teousness, which the Lord, the righteous judge, shall give me at that day: and not to me only, but unto all them also

that love his appearing. reposita est mihi iustitiae corona quam reddet mihi Domi-

nus in illa die iustus iudex non solum autem mihi sed et

4:9 Do thy diligence to come shortly unto me: his qui diligunt aduentum eius 4:9 festina uenire ad me 4:10 For Demas hath forsaken me, having loved this cito 4:10 Demas enim me dereliquit diligens hoc saeculum

present world, and is departed unto Thessalonica; Crescens to Galatia, Titus unto Dalmatia.

et abiit Thessalonicam Crescens in Galliam Titus in Dalma-

4:11 Only Luke is with me. Take Mark, and bring him tiam 4:11 Lucas est mecum solus Marcum adsume et adduc

with thee: for he is profitable to me for the ministry.

4:12 And Tychicus have I sent to Ephesus. tecum est enim mihi utilis in ministerium 4:12 Tychicum

4:13 The cloke that I left at Troas with Carpus, when thou autem misi Ephesum 4:13 paenulam quam reliqui Troade comest, bring with thee, and the books, but especially the parchments.

apud Carpum ueniens adfers et libros maxime autem mem-

4:14 Alexander the coppersmith did me much evil: the branas 4:14 Alexander aerarius multa mala mihi ostendit

Lord reward him according to his works: 4:15 Of whom be thou ware also; for he hath greatly with- reddat ei Dominus secundum opera eius 4:15 quem et tu stood our words.

4:16 At my first answer no man stood with me, but all deuita ualde enim restitit uerbis nostris 4:16 in prima mea

men forsook me: I pray God that it may not be laid to their charge.

defensione nemo mihi adfuit sed omnes me dereliquerunt

4:17 Notwithstanding the Lord stood with me, and non illis reputetur 4:17 Dominus autem mihi adstitit et con-

strengthened me; that by me the preaching might be fully known, and that all the Gentiles might hear: and I was

delivered out of the mouth of the lion. fortauit me ut per me praedicatio impleatur et audiant omnes 4:18 And the Lord shall deliver me from every evil work, gentes et liberatus sum de ore leonis 4:18 liberabit me Do-

and will preserve me unto his heavenly kingdom: to whom be glory for ever and ever. Amen.

minus ab omni opere malo et saluum faciet in regnum suum

Epistola Pauli Ad Timotheum II

caeleste cui gloria in saecula saeculorum amen 4:19 saluta 4:19 Salute Prisca and Aquila, and the household of One-

siphorus.

Priscam et Aquilam et Onesifori domum 4:20 Erastus re- 4:20 Erastus abode at Corinth: but Trophimus have I left

at Miletum sick.

mansit Corinthi Trophimum autem reliqui infirmum Mileti 4:21 festina ante hiemem uenire salutat te Eubulus et Pudens 4:21 Do thy diligence to come before winter. Eubulus

greeteth thee, and Pudens, and Linus, and Claudia, and all the brethren.

et Linus et Claudia et fratres omnes 4:22 Dominus Iesus

4:22 The Lord Jesus Christ be with thy spirit. Grace be with you. Amen.

cum spiritu tuo gratia nobiscum amen

Biblia Sacra Vulgata

MDCCXX

Epistola Pauli Ad Titum

1:1 Paulus seruus Dei apostolus autem Iesu Christi secun- 1:1 Paul, a servant of God, and an apostle of Jesus Christ,

according to the faith of God’s elect, and the acknowledging of the truth which is after godliness;

dum fidem electorum Dei et agnitionem ueritatis quae secundum pietatem est 1:2 in spem uitae aeternae quam pro- 1:2 In hope of eternal life, which God, that cannot lie,

promised before the world began;

misit qui non mentitur Deus ante tempora saecularia 1:3 1:3 But hath in due times manifested his word through

preaching, which is committed unto me according to the commandment of God our Saviour;

manifestauit autem temporibus suis uerbum suum in praedicatione quae credita est mihi secundum praeceptum saluatoris nostri Dei 1:4 Tito dilecto filio secundum communem 1:4 To Titus, mine own son after the common faith: Grace,

mercy, and peace, from God the Father and the Lord Jesus Christ our Saviour.

fidem gratia et pax a Deo Patre et Christo Iesu saluatore nostro 1:5 huius rei gratia reliqui te Cretae ut ea quae desunt 1:5 For this cause left I thee in Crete, that thou shouldest

set in order the things that are wanting, and ordain elders in every city, as I had appointed thee:

corrigas et constituas per ciuitates presbyteros sicut ego tibi disposui 1:6 si quis sine crimine est unius uxoris uir filios 1:6 If any be blameless, the husband of one wife, having

faithful children not accused of riot or unruly.

habens fideles non in accusatione luxuriae aut non subditos 1:7 oportet enim episcopum sine crimine esse sicut Dei di- 1:7 For a bishop must be blameless, as the steward of

God; not selfwilled, not soon angry, not given to wine, no striker, not given to filthy lucre;

spensatorem non superbum non iracundum non uinolentum non percussorem non turpilucri cupidum 1:8 sed hospita- 1:8 But a lover of hospitality, a lover of good men, sober,

just, holy, temperate;

lem benignum sobrium iustum sanctum continentem 1:9 1:9 Holding fast the faithful word as he hath been taught, that he may be able by sound doctrine both to exhort and to convince the gainsayers.

amplectentem eum qui secundum doctrinam est fidelem sermonem ut potens sit et exhortari in doctrina sana et eos qui contradicunt arguere 1:10 sunt enim multi et inoboedientes 1:10 For there are many unruly and vain talkers and de-

ceivers, specially they of the circumcision:

uaniloqui et seductores maxime qui de circumcisione sunt 1:11 quos oportet redargui qui uniuersas domos subuertunt 1:11 Whose mouths must be stopped, who subvert whole

houses, teaching things which they ought not, for filthy lucre’s sake.

docentes quae non oportet turpis lucri gratia 1:12 dixit qui-

1:12 One of themselves, even a prophet of their own, said, The Cretians are alway liars, evil beasts, slow bellies.

dam ex illis proprius ipsorum propheta Cretenses semper mendaces malae bestiae uentres pigri 1:13 testimonium hoc 1:13 This witness is true. Wherefore rebuke them sharply,

that they may be sound in the faith;

uerum est quam ob causam increpa illos dure ut sani sint in fide 1:14 non intendentes iudaicis fabulis et mandatis homi- 1:14 Not giving heed to Jewish fables, and commandments

of men, that turn from the truth.

num auersantium se a ueritate 1:15 omnia munda mundis 1:15 Unto the pure all things are pure: but unto them that are defiled and unbelieving is nothing pure; but even their mind and conscience is defiled.

coinquinatis autem et infidelibus nihil mundum sed inquinatae sunt eorum et mens et conscientia 1:16 confitentur se 1:16 They profess that they know God; but in works they

deny him, being abominable, and disobedient, and unto every good work reprobate.

nosse Deum factis autem negant cum sunt abominati et incredibiles et ad omne opus bonum reprobi

2:1 tu autem loquere quae decet sanam doctrinam 2:2 se- 2:1 But speak thou the things which become sound doc-

trine:

2:2 That the aged men be sober, grave, temperate, sound in faith, in charity, in patience.

nes ut sobrii sint pudici prudentes sani fide dilectione pa-

Biblia Sacra Vulgata

2:3 The aged women likewise, that they be in behaviour as tientia 2:3 anus similiter in habitu sancto non criminatri-

becometh holiness, not false accusers, not given to much wine, teachers of good things;

ces non uino multo seruientes bene docentes 2:4 ut pru-

2:4 That they may teach the young women to be sober, to love their husbands, to love their children,

dentiam doceant adulescentulas ut uiros suos ament filios

2:5 To be discreet, chaste, keepers at home, good, obedi- diligant 2:5 prudentes castas domus curam habentes benig-

ent to their own husbands, that the word of God be not blasphemed.

nas subditas suis uiris ut non blasphemetur uerbum Dei 2:6

2:6 Young men likewise exhort to be sober minded.

2:7 In all things shewing thyself a pattern of good works: iuuenes similiter hortare ut sobrii sint 2:7 in omnibus te ip-

in doctrine shewing uncorruptness, gravity, sincerity,

sum praebe exemplum bonorum operum in doctrina integ-

2:8 Sound speech, that cannot be condemned; that he that ritatem grauitatem 2:8 uerbum sanum inreprehensibilem ut

is of the contrary part may be ashamed, having no evil thing to say of you.

is qui ex aduerso est uereatur nihil habens malum dicere de

2:9 Exhort servants to be obedient unto their own masters, nobis 2:9 seruos dominis suis subditos esse in omnibus pla-

and to please them well in all things; not answering again; 2:10 Not purloining, but shewing all good fidelity; that centes non contradicentes 2:10 non sed in omnibus fidem they may adorn the doctrine of God our Saviour in all things.

bonam ostendentes ut doctrinam salutaris nostri Dei ornent

2:11 For the grace of God that bringeth salvation hath ap- in omnibus 2:11 apparuit enim gratia Dei salutaris omnipeared to all men,

2:12 Teaching us that, denying ungodliness and worldly bus hominibus 2:12 erudiens nos ut abnegantes impietatem

lusts, we should live soberly, righteously, and godly, in this present world;

et saecularia desideria sobrie et iuste et pie uiuamus in hoc

2:13 Looking for that blessed hope, and the glorious ap- saeculo 2:13 expectantes beatam spem et aduentum gloriae

pearing of the great God and our Saviour Jesus Christ;

2:14 Who gave himself for us, that he might redeem us magni Dei et saluatoris nostri Iesu Christi 2:14 qui dedit

from all iniquity, and purify unto himself a peculiar people, zealous of good works.

semet ipsum pro nobis ut nos rediieret ab omni iniquitate et mundaret sibi populum acceptabilem sectatorem bonorum

2:15 These things speak, and exhort, and rebuke with all operum 2:15 haec loquere et exhortare et argue cum omni authority. Let no man despise thee.

imperio nemo te contemnat

3:1 Put them in mind to be subject to principalities and 3:1 admone illos principibus et potestatibus subditos esse

powers, to obey magistrates, to be ready to every good work,

dicto oboedire ad omne opus bonum paratos esse 3:2 nemi-

3:2 To speak evil of no man, to be no brawlers, but gentle, shewing all meekness unto all men.

nem blasphemare non litigiosos esse modestos omnem os-

3:3 For we ourselves also were sometimes foolish, disobe- tendentes mansuetudinem ad omnes homines 3:3 eramus dient, deceived, serving divers lusts and pleasures, living in malice and envy, hateful, and hating one another.

enim et nos aliquando insipientes increduli errantes seruientes desideriis et uoluptatibus uariis in malitia et inuidia agen-

3:4 But after that the kindness and love of God our Saviour tes odibiles odientes inuicem 3:4 cum autem benignitas et

toward man appeared,

3:5 Not by works of righteousness which we have done, humanitas apparuit saluatoris nostri Dei 3:5 non ex operi-

but according to his mercy he saved us, by the washing of regeneration, and renewing of the Holy Ghost;

bus iustitiae quae fecimus nos sed secundum suam misericordiam saluos nos fecit per lauacrum regenerationis et re-

3:6 Which he shed on us abundantly through Jesus Christ nouationis Spiritus Sancti 3:6 quem effudit in nos abunde

our Saviour;

3:7 That being justified by his grace, we should be made per Iesum Christum saluatorem nostrum 3:7 ut iustificati

heirs according to the hope of eternal life.

Epistola Pauli Ad Titum

gratia ipsius heredes simus secundum spem uitae aeternae

3:8 fidelis sermo est et de his uolo te confirmare ut curent 3:8 This is a faithful saying, and these things I will that thou affirm constantly, that they which have believed in God might be careful to maintain good works. These

bonis operibus praeesse qui credunt Deo haec sunt bona et things are good and profitable unto men. utilia hominibus 3:9 stultas autem quaestiones et genealo- 3:9 But avoid foolish questions, and genealogies, and con-

tentions, and strivings about the law; for they are unprofitable and vain.

gias et contentiones et pugnas legis deuita sunt enim inutiles et uanae 3:10 hereticum hominem post unam et secun- 3:10 A man that is an heretick after the first and second

admonition reject;

dam correptionem deuita 3:11 sciens quia subuersus est qui 3:11 Knowing that he that is such is subverted, and sin-

neth, being condemned of himself.

eiusmodi est et delinquit proprio iudicio condemnatus 3:12 3:12 When I shall send Artemas unto thee, or Tychicus,

be diligent to come unto me to Nicopolis: for I have determined there to winter.

cum misero ad te Arteman aut Tychicum festina ad me uenire Nicopolim ibi enim statui hiemare 3:13 Zenan legis 3:13 Bring Zenas the lawyer and Apollos on their journey

diligently, that nothing be wanting unto them.

peritum et Apollo sollicite praemitte ut nihil illis desit 3:14 3:14 And let our’s also learn to maintain good works for

necessary uses, that they be not unfruitful.

discant autem et nostri bonis operibus praeesse ad usus necessarios ut non sint infructuosi 3:15 salutant te qui mecum 3:15 All that are with me salute thee. Greet them that love

us in the faith. Grace be with you all. Amen.

sunt omnes saluta qui nos amant in fide gratia Dei cum omnibus uobis amen

Biblia Sacra Vulgata

MDCCXXIV

Epistola Pauli Ad Philemonem

1:1 Paulus uinctus Iesu Christi et Timotheus frater Phile- 1:1 Paul, a prisoner of Jesus Christ, and Timothy our

brother, unto Philemon our dearly beloved, and fellowlabourer,

moni dilecto et adiutori nostro 1:2 et Appiae sorori et Ar-

1:2 And to our beloved Apphia, and Archippus our fellowsoldier, and to the church in thy house:

chippo commilitoni nostro et ecclesiae quae in domo tua est 1:3 gratia uobis et pax a Deo Patre nostro et Domino Iesu 1:3 Grace to you, and peace, from God our Father and the

Lord Jesus Christ.

Christo 1:4 gratias ago Deo meo semper memoriam tui fa- 1:4 I thank my God, making mention of thee always in my

prayers,

ciens in orationibus meis 1:5 audiens caritatem tuam et fi- 1:5 Hearing of thy love and faith, which thou hast toward

the Lord Jesus, and toward all saints;

dem quam habes in Domino Iesu et in omnes sanctos 1:6 ut 1:6 That the communication of thy faith may become ef-

fectual by the acknowledging of every good thing which is in you in Christ Jesus.

communicatio fidei tuae euidens fiat in agnitione omnis boni in nobis in Christo Iesu 1:7 gaudium enim magnum habui 1:7 For we have great joy and consolation in thy love,

because the bowels of the saints are refreshed by thee, brother.

et consolationem in caritate tua quia uiscera sanctorum requieuerunt per te frater 1:8 propter quod multam fiduciam 1:8 Wherefore, though I might be much bold in Christ to

enjoin thee that which is convenient,

habentes in Christo Iesu imperandi tibi quod ad rem pertinet 1:9 propter caritatem magis obsecro cum sis talis ut Paulus 1:9 Yet for love’s sake I rather beseech thee, being such

an one as Paul the aged, and now also a prisoner of Jesus Christ.

senex nunc autem et uinctus Iesu Christi 1:10 obsecro te

1:10 I beseech thee for my son Onesimus, whom I have begotten in my bonds:

de meo filio quem genui in uinculis Onesimo 1:11 qui tibi 1:11 Which in time past was to thee unprofitable, but now

profitable to thee and to me:

aliquando inutilis fuit nunc autem et tibi et mihi utilis 1:12 1:12 Whom I have sent again: thou therefore receive him,

that is, mine own bowels:

quem remisi tu autem illum id est mea uiscera suscipe 1:13 1:13 Whom I would have retained with me, that in thy

stead he might have ministered unto me in the bonds of the gospel:

quem ego uolueram mecum detinere ut pro te mihi ministraret in uinculis euangelii 1:14 sine consilio autem tuo ni- 1:14 But without thy mind would I do nothing; that thy

benefit should not be as it were of necessity, but willingly.

hil uolui facere uti ne uelut ex necessitate bonum tuum esset sed uoluntarium 1:15 forsitan enim ideo discessit ad horam 1:15 For perhaps he therefore departed for a season, that

thou shouldest receive him for ever;

a te ut aeternum illum recipere 1:16 iam non ut seruum sed 1:16 Not now as a servant, but above a servant, a brother

beloved, specially to me, but how much more unto thee, both in the flesh, and in the Lord?

plus seruo carissimum fratrem maxime mihi quanto autem

magis tibi et in carne et in Domino 1:17 si ergo habes me 1:17 If thou count me therefore a partner, receive him as

myself.

socium suscipe illum sicut me 1:18 si autem aliquid nocuit 1:18 If he hath wronged thee, or oweth thee ought, put that

on mine account;

tibi aut debet hoc mihi inputa 1:19 ego Paulus scripsi mea 1:19 I Paul have written it with mine own hand, I will re-

pay it: albeit I do not say to thee how thou owest unto me even thine own self besides.

manu ego reddam ut non dicam tibi quod et te ipsum mihi debes 1:20 ita frater ego te fruar in Domino refice uiscera 1:20 Yea, brother, let me have joy of thee in the Lord:

refresh my bowels in the Lord.

mea in Domino 1:21 confidens oboedientia tua scripsi tibi 1:21 Having confidence in thy obedience I wrote unto

thee, knowing that thou wilt also do more than I say.

sciens quoniam et super id quod dico facies 1:22 simul au- 1:22 But withal prepare me also a lodging: for I trust that

through your prayers I shall be given unto you.

tem et para mihi hospitium nam spero per orationes uestras donari me uobis 1:23 salutat te Epaphras concaptiuus meus 1:23 There salute thee Epaphras, my fellowprisoner in

Christ Jesus;

in Christo Iesu 1:24 Marcus Aristarchus Demas Lucas ad- 1:24 Marcus, Aristarchus, Demas, Lucas, my fel-

lowlabourers.

Biblia Sacra Vulgata

1:25 The grace of our Lord Jesus Christ be with your spirit. iutores mei 1:25 gratia Domini nostri Iesu Christi cum spi-

Amen.

ritu uestro amen

Epistola Pauli Ad Hebraeos

1:1 multifariam et multis modis olim Deus loquens patri- 1:1 God, who at sundry times and in divers manners spake

in time past unto the fathers by the prophets,

bus in prophetis 1:2 nouissime diebus istis locutus est no- 1:2 Hath in these last days spoken unto us by his Son,

whom he hath appointed heir of all things, by whom also he made the worlds;

bis in Filio quem constituit heredem uniuersorum per quem fecit et saecula 1:3 qui cum sit splendor gloriae et figura 1:3 Who being the brightness of his glory, and the express

image of his person, and upholding all things by the word of his power, when he had by himself purged our sins, sat

substantiae eius portansque omnia uerbo uirtutis suae pur- down on the right hand of the Majesty on high: gationem peccatorum faciens sedit ad dexteram Maiestatis in excelsis 1:4 tanto melior angelis effectus quanto diffe- 1:4 Being made so much better than the angels, as he hath

by inheritance obtained a more excellent name than they.

rentius prae illis nomen hereditauit 1:5 cui enim dixit ali- 1:5 For unto which of the angels said he at any time, Thou

art my Son, this day have I begotten thee? And again, I will be to him a Father, and he shall be to me a Son?

quando angelorum Filius meus es tu ego hodie genui te et rursum ego ero illi in Patrem et ipse erit mihi in Filium 1:6 1:6 And again, when he bringeth in the firstbegotten into

the world, he saith, And let all the angels of God worship him.

et cum iterum introducit primogenitum in orbem terrae dicit et adorent eum omnes angeli Dei 1:7 et ad angelos quidem 1:7 And of the angels he saith, Who maketh his angels

spirits, and his ministers a flame of fire.

dicit qui facit angelos suos spiritus et ministros suos flammam ignis 1:8 ad Filium autem thronus tuus Deus in sae- 1:8 But unto the Son he saith, Thy throne, O God, is for

ever and ever: a sceptre of righteousness is the sceptre of thy kingdom.

culum saeculi et uirga aequitatis uirga regni tui 1:9 dilexisti

1:9 Thou hast loved righteousness, and hated iniquity; therefore God, even thy God, hath anointed thee with the oil of gladness above thy fellows.

iustitiam et odisti iniquitatem propterea unxit te Deus Deus tuus oleo exultationis prae participibus tuis 1:10 et tu in 1:10 And, Thou, Lord, in the beginning hast laid the foun-

dation of the earth; and the heavens are the works of thine hands:

principio Domine terram fundasti et opera manuum tuarum sunt caeli 1:11 ipsi peribunt tu autem permanebis et omnes 1:11 They shall perish; but thou remainest; and they all

shall wax old as doth a garment;

ut uestimentum ueterescent 1:12 et uelut amictum inuolues 1:12 And as a vesture shalt thou fold them up, and they

shall be changed: but thou art the same, and thy years shall not fail.

eos et mutabuntur tu autem idem es et anni tui non deficient 1:13 ad quem autem angelorum dixit aliquando sede a dext- 1:13 But to which of the angels said he at any time, Sit on

my right hand, until I make thine enemies thy footstool?

ris meis quoadusque ponam inimicos tuos scabillum pedum

tuorum 1:14 nonne omnes sunt administratorii spiritus in 1:14 Are they not all ministering spirits, sent forth to min-

ister for them who shall be heirs of salvation?

ministerium missi propter eos qui hereditatem capient salutis

2:1 propterea abundantius oportet obseruare nos ea quae 2:1 Therefore we ought to give the more earnest heed to

the things which we have heard, lest at any time we should let them slip.

audiuimus ne forte pereffluamus 2:2 si enim qui per ange-

2:2 For if the word spoken by angels was stedfast, and every transgression and disobedience received a just recompence of reward;

los dictus est sermo factus est firmus et omnis praeuaricatio et inoboedientia accepit iustam mercedis retributionem 2:3 quomodo nos effugiemus si tantam neglexerimus salu- 2:3 How shall we escape, if we neglect so great salvation;

which at the first began to be spoken by the Lord, and was confirmed unto us by them that heard him;

tem quae cum initium accepisset enarrari per Dominum ab eis qui audierunt in nos confirmata est 2:4 contestante Deo 2:4 God also bearing them witness, both with signs and

wonders, and with divers miracles, and gifts of the Holy Ghost, according to his own will?

Biblia Sacra Vulgata

signis et portentis et uariis uirtutibus et Spiritus Sancti distri-

2:5 For unto the angels hath he not put in subjection the butionibus secundum suam uoluntatem 2:5 non enim ange-

world to come, whereof we speak.

2:6 But one in a certain place testified, saying, What is lis subiecit orbem terrae futurum de quo loquimur 2:6 te-

man, that thou art mindful of him? or the son of man that thou visitest him?

status est autem in quodam loco quis dicens quid est homo quod memor es eius aut filius hominis quoniam uisitas eum

2:7 Thou madest him a little lower than the angels; thou 2:7 minuisti eum paulo minus ab angelis gloria et honore

crownedst him with glory and honour, and didst set him over the works of thy hands:

coronasti eum et constituisti eum super opera manuum tua-

2:8 Thou hast put all things in subjection under his feet. rum 2:8 omnia subiecisti sub pedibus eius in eo enim quod For in that he put all in subjection under him, he left nothing that is not put under him. But now we see not yet all

things put under him. ei omnia subiecit nihil dimisit non subiectum ei nunc au-

2:9 But we see Jesus, who was made a little lower than the tem necdum uidemus omnia subiecta ei 2:9 eum autem qui angels for the suffering of death, crowned with glory and honour; that he by the grace of God should taste death for

every man. modico quam angeli minoratus est uidemus Iesum propter

passionem mortis gloria et honore coronatum ut gratia Dei

2:10 For it became him, for whom are all things, and by pro omnibus gustaret mortem 2:10 decebat enim eum pro-

whom are all things, in bringing many sons unto glory, to make the captain of their salvation perfect through suffer-

ings. pter quem omnia et per quem omnia qui multos filios in glo-

riam adduxerat auctorem salutis eorum per passiones con-

2:11 For both he that sanctifieth and they who are sancti- summare 2:11 qui enim sanctificat et qui sanctificantur ex

fied are all of one: for which cause he is not ashamed to call them brethren,

uno omnes propter quam causam non confunditur fratres eos

2:12 Saying, I will declare thy name unto my brethren, in uocare dicens 2:12 nuntiabo nomen tuum fratribus meis in

the midst of the church will I sing praise unto thee.

2:13 And again, I will put my trust in him. And again, medio ecclesiae laudabo te 2:13 et iterum ego ero fidens in

Behold I and the children which God hath given me.

eum et iterum ecce ego et pueri mei quos mihi dedit Deus

2:14 Forasmuch then as the children are partakers of flesh 2:14 quia ergo pueri communicauerunt sanguini et carni et

and blood, he also himself likewise took part of the same; that through death he might destroy him that had the power

of death, that is, the devil; ipse similiter participauit hisdem ut per mortem destrueret

2:15 And deliver them who through fear of death were all eum qui habebat mortis imperium id est diabolum 2:15 et their lifetime subject to bondage.

liberaret eos qui timore mortis per totam uitam obnoxii erant

2:16 For verily he took not on him the nature of angels; seruituti 2:16 nusquam enim angelos adprehendit sed se-

but he took on him the seed of Abraham.

2:17 Wherefore in all things it behoved him to be made men Abrahae adprehendit 2:17 unde debuit per omnia frat-

like unto his brethren, that he might be a merciful and faithful high priest in things pertaining to God, to make

reconciliation for the sins of the people. ribus similare ut misericors fieret et fidelis pontifex ad Deum 2:18 For in that he himself hath suffered being tempted, he ut repropitiaret delicta populi 2:18 in eo enim in quo passus

is able to succour them that are tempted.

est ipse temptatus potens est eis qui temptantur auxiliari

3:1 Wherefore, holy brethren, partakers of the heavenly 3:1 unde fratres sancti uocationis caelestis participes con-

calling, consider the Apostle and High Priest of our profession, Christ Jesus;

siderate apostolum et pontificem confessionis nostrae Iesum

3:2 Who was faithful to him that appointed him, as also 3:2 qui fidelis est ei qui fecit illum sicut et Moses in omni

Moses was faithful in all his house.

3:3 For this man was counted worthy of more glory than domo illius 3:3 amplioris enim gloriae iste prae Mose di-

Moses, inasmuch as he who hath builded the house hath more honour than the house.

Epistola Pauli Ad Hebraeos

gnus habitus est quanto ampliorem honorem habet domus qui fabricauit illam 3:4 omnis namque domus fabricatur ab 3:4 For every house is builded by some man; but he that

built all things is God.

aliquo qui autem omnia creauit Deus 3:5 et Moses quidem 3:5 And Moses verily was faithful in all his house, as a

servant, for a testimony of those things which were to be spoken after;

fidelis erat in tota domo eius tamquam famulus in testimonium eorum quae dicenda erant 3:6 Christus uero tamquam 3:6 But Christ as a son over his own house; whose house

are we, if we hold fast the confidence and the rejoicing of the hope firm unto the end.

filius in domo sua quae domus sumus nos si fiduciam et glo-

riam spei usque ad finem firmam retineamus 3:7 quaprop- 3:7 Wherefore (as the Holy Ghost saith, To day if ye will

hear his voice,

ter sicut dicit Spiritus Sanctus hodie si uocem eius audieritis 3:8 nolite obdurare corda uestra sicut in exacerbatione secun- 3:8 Harden not your hearts, as in the provocation, in the

day of temptation in the wilderness:

dum diem temptationis in deserto 3:9 ubi temptauerunt me 3:9 When your fathers tempted me, proved me, and saw

my works forty years.

patres uestri probauerunt et uiderunt opera mea 3:10 quad- 3:10 Wherefore I was grieved with that generation, and

said, They do alway err in their heart; and they have not known my ways.

raginta annos propter quod infensus fui generationi huic et dixi semper errant corde ipsi autem non cognouerunt uias meas 3:11 sicut iuraui in ira mea si introibunt in requiem 3:11 So I sware in my wrath, They shall not enter into my

rest.)

meam 3:12 uidete fratres ne forte sit in aliquo uestrum cor 3:12 Take heed, brethren, lest there be in any of you an

evil heart of unbelief, in departing from the living God.

malum incredulitatis discedendi a Deo uiuo 3:13 sed adhor- 3:13 But exhort one another daily, while it is called To

day; lest any of you be hardened through the deceitfulness of sin.

tamini uosmet ipsos per singulos dies donec hodie cognominatur ut non obduretur quis ex uobis fallacia peccati 3:14 3:14 For we are made partakers of Christ, if we hold the

beginning of our confidence stedfast unto the end;

participes enim Christi effecti sumus si tamen initium substantiae usque ad finem firmum retineamus 3:15 dum dicitur 3:15 While it is said, To day if ye will hear his voice,

harden not your hearts, as in the provocation.

hodie si uocem eius audieritis nolite obdurare corda uestra quemadmodum in illa exacerbatione 3:16 quidam enim au- 3:16 For some, when they had heard, did provoke: howbeit

not all that came out of Egypt by Moses.

dientes exacerbauerunt sed non uniuersi qui profecti sunt ab Aegypto per Mosen 3:17 quibus autem infensus est quad- 3:17 But with whom was he grieved forty years? was it

not with them that had sinned, whose carcases fell in the wilderness?

raginta annos nonne illis qui peccauerunt quorum cadauera prostrata sunt in deserto 3:18 quibus autem iurauit non in- 3:18 And to whom sware he that they should not enter into

his rest, but to them that believed not?

troire in requiem ipsius nisi illis qui increduli fuerunt 3:19 3:19 So we see that they could not enter in because of

unbelief.

et uidemus quia non potuerunt introire propter incredulitatem

4:1 timeamus ergo ne forte relicta pollicitatione introeundi 4:1 Let us therefore fear, lest, a promise being left us of

entering into his rest, any of you should seem to come short of it.

in requiem eius existimetur aliqui ex uobis deesse 4:2 ete-

4:2 For unto us was the gospel preached, as well as unto them: but the word preached did not profit them, not being mixed with faith in them that heard it.

nim et nobis nuntiatum est quemadmodum et illis sed non profuit illis sermo auditus non admixtis fidei ex his quae

Biblia Sacra Vulgata

4:3 For we which have believed do enter into rest, as he audierunt 4:3 ingrediemur enim in requiem qui credidimus

said, As I have sworn in my wrath, if they shall enter into my rest: although the works were finished from the foun-

dation of the world. quemadmodum dixit sicut iuraui in ira mea si introibunt in

requiem meam et quidem operibus ab institutione mundi fac-

4:4 For he spake in a certain place of the seventh day on tis 4:4 dixit enim quodam loco de die septima sic et rethis wise, And God did rest the seventh day from all his works.

quieuit Deus die septima ab omnibus operibus suis 4:5 et

4:5 And in this place again, If they shall enter into my rest.

4:6 Seeing therefore it remaineth that some must enter in isto rursum si introibunt in requiem meam 4:6 quoniam

therein, and they to whom it was first preached entered not in because of unbelief:

ergo superest quosdam introire in illam et hii quibus prioribus adnuntiatum est non introierunt propter incredulitatem

4:7 Again, he limiteth a certain day, saying in David, To 4:7 iterum terminat diem quendam hodie in Dauid dicendo day, after so long a time; as it is said, To day if ye will hear his voice, harden not your hearts.

post tantum temporis sicut supra dictum est hodie si uo-

4:8 For if Jesus had given them rest, then would he not cem eius audieritis nolite obdurare corda uestra 4:8 nam si

afterward have spoken of another day.

eis Iesus requiem praestitisset numquam de alio loqueretur

4:9 There remaineth therefore a rest to the people of God. posthac die 4:9 itaque relinquitur sabbatismus populo Dei

4:10 For he that is entered into his rest, he also hath ceased 4:10 qui enim ingressus est in requiem eius etiam ipse refrom his own works, as God did from his. 4:11 Let us labour therefore to enter into that rest, lest any quieuit ab operibus suis sicut a suis Deus 4:11 festinemus man fall after the same example of unbelief.

ergo ingredi in illam requiem ut ne in id ipsum quis inci-

4:12 For the word of God is quick, and powerful, and dat incredulitatis exemplum 4:12 uiuus est enim Dei sermo

sharper than any twoedged sword, piercing even to the dividing asunder of soul and spirit, and of the joints and marrow, and is a discerner of the thoughts and intents of the heart.

et efficax et penetrabilior omni gladio ancipiti et pertingens usque ad diuisionem animae ac spiritus conpagum quoque et medullarum et discretor cogitationum et intentionum cor-

4:13 Neither is there any creature that is not manifest in his dis 4:13 et non est ulla creatura inuisibilis in conspectu eius

sight: but all things are naked and opened unto the eyes of him with whom we have to do.

omnia autem nuda et aperta sunt oculis eius ad quem no-

4:14 Seeing then that we have a great high priest, that is bis sermo 4:14 habentes ergo pontificem magnum qui pepassed into the heavens, Jesus the Son of God, let us hold fast our profession.

netrauerit caelos Iesum Filium Dei teneamus confessionem

4:15 For we have not an high priest which cannot be 4:15 non enim habemus pontificem qui non possit conpati

touched with the feeling of our infirmities; but was in all points tempted like as we are, yet without sin.

infirmitatibus nostris temptatum autem per omnia pro simi-

4:16 Let us therefore come boldly unto the throne of grace, litudine absque peccato 4:16 adeamus ergo cum fiducia ad that we may obtain mercy, and find grace to help in time of need.

thronum gratiae ut misericordiam consequamur et gratiam inueniamus in auxilio oportuno

5:1 For every high priest taken from among men is or- 5:1 omnis namque pontifex ex hominibus adsumptus pro

dained for men in things pertaining to God, that he may offer both gifts and sacrifices for sins:

hominibus constituitur in his quae sunt ad Deum ut offerat

5:2 Who can have compassion on the ignorant, and on dona et sacrificia pro peccatis 5:2 qui condolere possit his

them that are out of the way; for that he himself also is compassed with infirmity.

qui ignorant et errant quoniam et ipse circumdatus est infir-

Epistola Pauli Ad Hebraeos

mitate 5:3 et propter eam debet quemadmodum et pro po- 5:3 And by reason hereof he ought, as for the people, so

also for himself, to offer for sins.

pulo ita etiam pro semet ipso offerre pro peccatis 5:4 nec 5:4 And no man taketh this honour unto himself, but he

that is called of God, as was Aaron.

quisquam sumit sibi honorem sed qui uocatur a Deo tamquam Aaron 5:5 sic et Christus non semet ipsum clarifica- 5:5 So also Christ glorified not himself to be made an high

priest; but he that said unto him, Thou art my Son, to day have I begotten thee.

uit ut pontifex fieret sed qui locutus est ad eum Filius meus

es tu ego hodie genui te 5:6 quemadmodum et in alio di- 5:6 As he saith also in another place, Thou art a priest for

ever after the order of Melchisedec.

cit tu es sacerdos in aeternum secundum ordinem Melchisedech 5:7 qui in diebus carnis suae preces supplicationesque 5:7 Who in the days of his flesh, when he had offered

up prayers and supplications with strong crying and tears unto him that was able to save him from death, and was

ad eum qui possit saluum illum a morte facere cum clamore heard in that he feared; ualido et lacrimis offerens et exauditus pro sua reuerentia 5:8 et quidem cum esset Filius didicit ex his quae passus 5:8 Though he were a Son, yet learned he obedience by

the things which he suffered;

est oboedientiam 5:9 et consummatus factus est omnibus 5:9 And being made perfect, he became the author of eter-

nal salvation unto all them that obey him;

obtemperantibus sibi causa salutis aeternae 5:10 appellatus 5:10 Called of God an high priest after the order of

Melchisedec.

a Deo pontifex iuxta ordinem Melchisedech 5:11 de quo 5:11 Of whom we have many things to say, and hard to be

uttered, seeing ye are dull of hearing.

grandis nobis sermo et ininterpretabilis ad dicendum quoniam inbecilles facti estis ad audiendum 5:12 etenim cum 5:12 For when for the time ye ought to be teachers, ye have

need that one teach you again which be the first principles of the oracles of God; and are become such as have need

deberetis magistri esse propter tempus rursum indigetis ut of milk, and not of strong meat. uos doceamini quae sint elementa exordii sermonum Dei et facti estis quibus lacte opus sit non solido cibo 5:13 om- 5:13 For every one that useth milk is unskilful in the word

of righteousness: for he is a babe.

nis enim qui lactis est particeps expers est sermonis iustitiae paruulus enim est 5:14 perfectorum autem est solidus cibus 5:14 But strong meat belongeth to them that are of full

age, even those who by reason of use have their senses exercised to discern both good and evil.

eorum qui pro consuetudine exercitatos habent sensus ad discretionem boni ac mali

6:1 quapropter intermittentes inchoationis Christi sermo- 6:1 Therefore leaving the principles of the doctrine of

Christ, let us go on unto perfection; not laying again the foundation of repentance from dead works, and of faith

nem ad perfectionem feramur non rursum iacientes funda- toward God, mentum paenitentiae ab operibus mortuis et fidei ad Deum 6:2 baptismatum doctrinae inpositionis quoque manuum ac 6:2 Of the doctrine of baptisms, and of laying on of hands,

and of resurrection of the dead, and of eternal judgment.

resurrectionis mortuorum et iudicii aeterni 6:3 et hoc facie- 6:3 And this will we do, if God permit. mus siquidem permiserit Deus 6:4 inpossibile est enim eos 6:4 For it is impossible for those who were once enlight-

ened, and have tasted of the heavenly gift, and were made partakers of the Holy Ghost,

qui semel sunt inluminati gustauerunt etiam donum caeleste et participes sunt facti Spiritus Sancti 6:5 gustauerunt nihi- 6:5 And have tasted the good word of God, and the powers

of the world to come,

lominus bonum Dei uerbum uirtutesque saeculi uenturi 6:6 6:6 If they shall fall away, to renew them again unto repen-

tance; seeing they crucify to themselves the Son of God afresh, and put him to an open shame.

et prolapsi sunt renouari rursus ad paenitentiam rursum cru-

Biblia Sacra Vulgata

6:7 For the earth which drinketh in the rain that cometh oft cifigentes sibimet ipsis Filium Dei et ostentui habentes 6:7

upon it, and bringeth forth herbs meet for them by whom it is dressed, receiveth blessing from God:

terra enim saepe uenientem super se bibens imbrem et generans herbam oportunam illis a quibus colitur accipit bene-

6:8 But that which beareth thorns and briers is rejected, dictionem a Deo 6:8 proferens autem spinas ac tribulos rep-

and is nigh unto cursing; whose end is to be burned.

roba est et maledicto proxima cuius consummatio in conbus-

6:9 But, beloved, we are persuaded better things of you, tionem 6:9 confidimus autem de uobis dilectissimi meliora

and things that accompany salvation, though we thus speak.

et uiciniora saluti tametsi ita loquimur 6:10 non enim in-

6:10 For God is not unrighteous to forget your work and labour of love, which ye have shewed toward his name, in that ye have ministered to the saints, and do minister.

iustus Deus ut obliuiscatur operis uestri et dilectionis quam ostendistis in nomine ipsius qui ministrastis sanctis et mini-

6:11 And we desire that every one of you do shew the same stratis 6:11 cupimus autem unumquemque uestrum eandem

diligence to the full assurance of hope unto the end:

ostentare sollicitudinem ad expletionem spei usque in fi-

6:12 That ye be not slothful, but followers of them who nem 6:12 ut non segnes efficiamini uerum imitatores eorum

through faith and patience inherit the promises.

6:13 For when God made promise to Abraham, because qui fide et patientia hereditabunt promissiones 6:13 Abra-

he could swear by no greater, he sware by himself,

hae namque promittens Deus quoniam neminem habuit per

6:14 Saying, Surely blessing I will bless thee, and multi- quem iuraret maiorem iurauit per semet ipsum 6:14 dicens plying I will multiply thee.

nisi benedicens benedicam te et multiplicans multiplicabo

6:15 And so, after he had patiently endured, he obtained te 6:15 et sic longanimiter ferens adeptus est repromissio-

the promise. 6:16 For men verily swear by the greater: and an oath for nem 6:16 homines enim per maiorem sui iurant et omnis confirmation is to them an end of all strife.

controuersiae eorum finis ad confirmationem est iuramen-

6:17 Wherein God, willing more abundantly to shew unto tum 6:17 in quo abundantius uolens Deus ostendere polthe heirs of promise the immutability of his counsel, confirmed it by an oath:

licitationis heredibus inmobilitatem consilii sui interposuit

6:18 That by two immutable things, in which it was im- iusiurandum 6:18 ut per duas res inmobiles quibus inpos-

possible for God to lie, we might have a strong consolation, who have fled for refuge to lay hold upon the hope

set before us: sibile est mentiri Deum fortissimum solacium habeamus qui 6:19 Which hope we have as an anchor of the soul, both confugimus ad tenendam propositam spem 6:19 quam sicut

sure and stedfast, and which entereth into that within the veil;

anchoram habemus animae tutam ac firmam et incedentem

6:20 Whither the forerunner is for us entered, even Je- usque in interiora uelaminis 6:20 ubi praecursor pro nobis

sus, made an high priest for ever after the order of Melchisedec.

introiit Iesus secundum ordinem Melchisedech pontifex factus in aeternum

7:1 For this Melchisedec, king of Salem, priest of the most 7:1 hic enim Melchisedech rex Salem sacerdos Dei summi high God, who met Abraham returning from the slaughter of the kings, and blessed him;

qui obuiauit Abrahae regresso a caede regum et benedixit

7:2 To whom also Abraham gave a tenth part of all; first ei 7:2 cui decimas omnium diuisit Abraham primum qui-

being by interpretation King of righteousness, and after that also King of Salem, which is, King of peace;

dem qui interpretatur rex iustitiae deinde autem et rex Salem

7:3 Without father, without mother, without descent, hav- quod est rex pacis 7:3 sine patre sine matre sine genealogia ing neither beginning of days, nor end of life; but made like unto the Son of God; abideth a priest continually.

Epistola Pauli Ad Hebraeos

neque initium dierum neque finem uitae habens adsimilatus autem Filio Dei manet sacerdos in perpetuum 7:4 in- 7:4 Now consider how great this man was, unto whom

even the patriarch Abraham gave the tenth of the spoils.

tuemini autem quantus sit hic cui et decimam dedit de praecipuis Abraham patriarcha 7:5 et quidem de filiis Leui sa- 7:5 And verily they that are of the sons of Levi, who re-

ceive the office of the priesthood, have a commandment to take tithes of the people according to the law, that is, of their brethren, though they come out of the loins of Abraham:

cerdotium accipientes mandatum habent decimas sumere a populo secundum legem id est a fratribus suis quamquam et ipsi exierunt de lumbis Abrahae 7:6 cuius autem generatio 7:6 But he whose descent is not counted from them re-

ceived tithes of Abraham, and blessed him that had the promises.

non adnumeratur in eis decimas sumpsit Abraham et hunc

qui habebat repromissiones benedixit 7:7 sine ulla autem 7:7 And without all contradiction the less is blessed of the

better.

contradictione quod minus est a meliore benedicitur 7:8 et 7:8 And here men that die receive tithes; but there he re-

ceiveth them, of whom it is witnessed that he liveth.

hic quidem decimas morientes homines accipiunt ibi autem contestatus quia uiuit 7:9 et ut ita dictum sit per Abraham 7:9 And as I may so say, Levi also, who receiveth tithes,

payed tithes in Abraham.

et Leui qui decimas accipit decimatus est 7:10 adhuc enim 7:10 For he was yet in the loins of his father, when

Melchisedec met him.

in lumbis patris erat quando obuiauit ei Melchisedech 7:11 7:11 If therefore perfection were by the Levitical priest-

hood, (for under it the people received the law,) what further need was there that another priest should rise after the order of Melchisedec, and not be called after the order of Aaron?

si ergo consummatio per sacerdotium leuiticum erat populus enim sub ipso legem accepit quid adhuc necessarium secundum ordinem Melchisedech alium surgere sacerdotem et non secundum ordinem Aaron dici 7:12 translato enim sa- 7:12 For the priesthood being changed, there is made of

necessity a change also of the law.

cerdotio necesse est ut et legis translatio fiat 7:13 in quo 7:13 For he of whom these things are spoken pertaineth

to another tribe, of which no man gave attendance at the altar.

enim haec dicuntur de alia tribu est de qua nullus altario praesto fuit 7:14 manifestum enim quod ex Iuda ortus sit 7:14 For it is evident that our Lord sprang out of Juda; of

which tribe Moses spake nothing concerning priesthood.

Dominus noster in qua tribu nihil de sacerdotibus Moses

locutus est 7:15 et amplius adhuc manifestum est si se- 7:15 And it is yet far more evident: for that after the simil-

itude of Melchisedec there ariseth another priest,

cundum similitudinem Melchisedech exsurgit alius sacerdos 7:16 qui non secundum legem mandati carnalis factus 7:16 Who is made, not after the law of a carnal command-

ment, but after the power of an endless life.

est sed secundum uirtutem uitae insolubilis 7:17 contesta- 7:17 For he testifieth, Thou art a priest for ever after the

order of Melchisedec.

tur enim quoniam tu es sacerdos in aeternum secundum ordinem Melchisedech 7:18 reprobatio quidem fit praecedentis 7:18 For there is verily a disannulling of the command-

ment going before for the weakness and unprofitableness thereof.

mandati propter infirmitatem eius et inutilitatem 7:19 ni-

7:19 For the law made nothing perfect, but the bringing in of a better hope did; by the which we draw nigh unto God.

hil enim ad perfectum adduxit lex introductio uero melioris spei per quam proximamus ad Deum 7:20 et quantum est 7:20 And inasmuch as not without an oath he was made

priest:

non sine iureiurando alii quidem sine iureiurando sacerdotes facti sunt 7:21 hic autem cum iureiurando per eum qui dixit 7:21 (For those priests were made without an oath; but this

with an oath by him that said unto him, The Lord sware and will not repent, Thou art a priest for ever after the order of Melchisedec:)

Biblia Sacra Vulgata

ad illum iurauit Dominus et non paenitebit tu es sacerdos in

7:22 By so much was Jesus made a surety of a better tes- aeternum 7:22 in tantum melioris testamenti sponsor fac-

tament.

7:23 And they truly were many priests, because they were tus est Iesus 7:23 et alii quidem plures facti sunt sacerdotes

not suffered to continue by reason of death:

7:24 But this man, because he continueth ever, hath an idcirco quod morte prohiberentur permanere 7:24 hic au-

unchangeable priesthood.

tem eo quod maneat in aeternum sempiternum habet sacer-

7:25 Wherefore he is able also to save them to the utter- dotium 7:25 unde et saluare in perpetuo potest accedentes

most that come unto God by him, seeing he ever liveth to make intercession for them.

per semet ipsum ad Deum semper uiuens ad interpellandum

7:26 For such an high priest became us, who is holy, harm- pro eis 7:26 talis enim decebat ut nobis esset pontifex sanc-

less, undefiled, separate from sinners, and made higher than the heavens;

tus innocens inpollutus segregatus a peccatoribus et excel-

7:27 Who needeth not daily, as those high priests, to offer sior caelis factus 7:27 qui non habet cotidie necessitatem

up sacrifice, first for his own sins, and then for the people’s: for this he did once, when he offered up himself.

quemadmodum sacerdotes prius pro suis delictis hostias offerre deinde pro populi hoc enim fecit semel se offerendo

7:28 For the law maketh men high priests which have in- 7:28 lex enim homines constituit sacerdotes infirmitatem ha-

firmity; but the word of the oath, which was since the law, maketh the Son, who is consecrated for evermore.

bentes sermo autem iurisiurandi qui post legem est Filium in aeternum perfectum

8:1 Now of the things which we have spoken this is the 8:1 capitulum autem super ea quae dicuntur talem habe-

sum: We have such an high priest, who is set on the right hand of the throne of the Majesty in the heavens;

mus pontificem qui consedit in dextera sedis Magnitudinis

8:2 A minister of the sanctuary, and of the true tabernacle, in caelis 8:2 sanctorum minister et tabernaculi ueri quod which the Lord pitched, and not man. 8:3 For every high priest is ordained to offer gifts and fixit Dominus et non homo 8:3 omnis enim pontifex ad ofsacrifices: wherefore it is of necessity that this man have somewhat also to offer.

ferenda munera et hostias constituitur unde necesse est et

8:4 For if he were on earth, he should not be a priest, see- hunc habere aliquid quod offerat 8:4 si ergo esset super ter-

ing that there are priests that offer gifts according to the law:

ram nec esset sacerdos cum essent qui offerrent secundum

8:5 Who serve unto the example and shadow of heavenly legem munera 8:5 qui exemplari et umbrae deseruiunt cae-

things, as Moses was admonished of God when he was about to make the tabernacle: for, See, saith he, that thou make all things according to the pattern shewed to thee in the mount.

lestium sicut responsum est Mosi cum consummaret tabernaculum uide inquit omnia facito secundum exemplar quod

8:6 But now hath he obtained a more excellent ministry, tibi ostensum est in monte 8:6 nunc autem melius sortitus

by how much also he is the mediator of a better covenant, which was established upon better promises.

est ministerium quanto et melioris testamenti mediator est

8:7 For if that first covenant had been faultless, then should quod in melioribus repromissionibus sanctum est 8:7 nam no place have been sought for the second.

si illud prius culpa uacasset non utique secundi locus inqui-

8:8 For finding fault with them, he saith, Behold, the days reretur 8:8 uituperans enim eos dicit ecce dies ueniunt di-

come, saith the Lord, when I will make a new covenant with the house of Israel and with the house of Judah:

cit Dominus et consummabo super domum Israhel et super

8:9 Not according to the covenant that I made with their domum Iuda testamentum nouum 8:9 non secundum testafathers in the day when I took them by the hand to lead them out of the land of Egypt; because they continued not

in my covenant, and I regarded them not, saith the Lord. mentum quod feci patribus eorum in die qua adprehendi ma-

Epistola Pauli Ad Hebraeos

num illorum ut educerem illos de terra Aegypti quoniam ipsi non permanserunt in testamento meo et ego neglexi eos dicit Dominus 8:10 quia hoc testamentum quod disponam do- 8:10 For this is the covenant that I will make with the

house of Israel after those days, saith the Lord; I will put my laws into their mind, and write them in their hearts: and I will be to them a God, and they shall be to me a people:

mui Israhel post dies illos dicit Dominus dando leges meas in mentem eorum et in corde eorum superscribam eas et ero eis in Deum et ipsi erunt mihi in populum 8:11 et non do- 8:11 And they shall not teach every man his neighbour,

and every man his brother, saying, Know the Lord: for all shall know me, from the least to the greatest.

cebit unusquisque proximum suum et unusquisque fratrem suum dicens cognosce Dominum quoniam omnes scient me

a minore usque ad maiorem eorum 8:12 quia propitius ero 8:12 For I will be merciful to their unrighteousness, and

their sins and their iniquities will I remember no more.

iniquitatibus eorum et peccatorum illorum iam non memorabor 8:13 dicendo autem nouum ueterauit prius quod autem 8:13 In that he saith, A new covenant, he hath made the

first old. Now that which decayeth and waxeth old is ready to vanish away.

antiquatur et senescit prope interitum est

9:1 habuit quidem et prius iustificationes culturae et sanc- 9:1 Then verily the first covenant had also ordinances of

divine service, and a worldly sanctuary.

tum saeculare 9:2 tabernaculum enim factum est primum 9:2 For there was a tabernacle made; the first, wherein was the candlestick, and the table, and the shewbread; which is called the sanctuary.

in quo inerant candelabra et mensa et propositio panum quae dicitur sancta 9:3 post uelamentum autem secundum taber- 9:3 And after the second veil, the tabernacle which is

called the Holiest of all;

naculum quod dicitur sancta sanctorum 9:4 aureum habens 9:4 Which had the golden censer, and the ark of the

covenant overlaid round about with gold, wherein was the golden pot that had manna, and Aaron’s rod that budded,

turibulum et arcam testamenti circumtectam ex omni parte and the tables of the covenant; auro in qua urna aurea habens manna et uirga Aaron quae fronduerat et tabulae testamenti 9:5 superque eam cherubin 9:5 And over it the cherubims of glory shadowing the mer-

cyseat; of which we cannot now speak particularly.

gloriae obumbrantia propitiatorium de quibus non est modo dicendum per singula 9:6 his uero ita conpositis in priori 9:6 Now when these things were thus ordained, the priests

went always into the first tabernacle, accomplishing the service of God.

quidem tabernaculo semper introibant sacerdotes sacrificio-

rum officia consummantes 9:7 in secundo autem semel in 9:7 But into the second went the high priest alone once every year, not without blood, which he offered for himself, and for the errors of the people:

anno solus pontifex non sine sanguine quem offert pro sua et populi ignorantia 9:8 hoc significante Spiritu Sancto non- 9:8 The Holy Ghost this signifying, that the way into the

holiest of all was not yet made manifest, while as the first tabernacle was yet standing:

dum propalatam esse sanctorum uiam adhuc priore tabernaculo habente statum 9:9 quae parabola est temporis instan- 9:9 Which was a figure for the time then present, in which

were offered both gifts and sacrifices, that could not make him that did the service perfect, as pertaining to the con-

tis iuxta quam munera et hostiae offeruntur quae non pos- science; sunt iuxta conscientiam perfectum facere seruientem 9:10 9:10 Which stood only in meats and drinks, and divers

washings, and carnal ordinances, imposed on them until the time of reformation.

solummodo in cibis et in potibus et uariis baptismis et iustitiis carnis usque ad tempus correctionis inpositis 9:11 9:11 But Christ being come an high priest of good things

to come, by a greater and more perfect tabernacle, not made with hands, that is to say, not of this building;

Christus autem adsistens pontifex futurorum bonorum per

Biblia Sacra Vulgata

amplius et perfectius tabernaculum non manufactum id est

9:12 Neither by the blood of goats and calves, but by his non huius creationis 9:12 neque per sanguinem hircorum

own blood he entered in once into the holy place, having obtained eternal redemption for us.

et uitulorum sed per proprium sanguinem introiuit semel in

9:13 For if the blood of bulls and of goats, and the ashes sancta aeterna redeeptione inuenta 9:13 si enim sanguis hir-

of an heifer sprinkling the unclean, sanctifieth to the purifying of the flesh:

corum et taurorum et cinis uitulae aspersus inquinatos sanc-

9:14 How much more shall the blood of Christ, who tificat ad emundationem carnis 9:14 quanto magis sanguis

through the eternal Spirit offered himself without spot to God, purge your conscience from dead works to serve the

living God? Christi qui per Spiritum Sanctum semet ipsum obtulit inma-

culatum Deo emundabit conscientiam uestram ab operibus

9:15 And for this cause he is the mediator of the new mortuis ad seruiendum Deo uiuenti 9:15 et ideo noui tes-

testament, that by means of death, for the redemption of the transgressions that were under the first testament, they which are called might receive the promise of eternal inheritance.

tamenti mediator est ut morte intercedente in redeeptionem earum praeuaricationum quae erant sub priore testamento repromissionem accipiant qui uocati sunt aeternae heredita-

9:16 For where a testament is, there must also of necessity tis 9:16 ubi enim testamentum mors necesse est intercedat be the death of the testator. 9:17 For a testament is of force after men are dead: other- testatoris 9:17 testamentum enim in mortuis confirmatum wise it is of no strength at all while the testator liveth. 9:18 Whereupon neither the first testament was dedicated est alioquin nondum ualet dum uiuit qui testatus est 9:18 without blood.

9:19 For when Moses had spoken every precept to all the unde ne primum quidem sine sanguine dedicatum est 9:19

people according to the law, he took the blood of calves and of goats, with water, and scarlet wool, and hyssop,

and sprinkled both the book, and all the people, lecto enim omni mandato legis a Mose uniuerso populo accipiens sanguinem uitulorum et hircorum cum aqua et lana coccinea et hysopo ipsum quoque librum et omnem popu-

9:20 Saying, This is the blood of the testament which God lum aspersit 9:20 dicens hic sanguis testamenti quod man-

hath enjoined unto you. 9:21 Moreover he sprinkled with blood both the taberna- dauit ad uos Deus 9:21 etiam tabernaculum et omnia uasa cle, and all the vessels of the ministry.

9:22 And almost all things are by the law purged with ministerii sanguine similiter aspersit 9:22 et omnia paene

blood; and without shedding of blood is no remission.

in sanguine mundantur secundum legem et sine sanguinis

9:23 It was therefore necessary that the patterns of things fusione non fit remissio 9:23 necesse est ergo exemplaria in the heavens should be purified with these; but the heavenly things themselves with better sacrifices than these.

quidem caelestium his mundari ipsa autem caelestia melio-

9:24 For Christ is not entered into the holy places made ribus hostiis quam istis 9:24 non enim in manufactis sanc-

with hands, which are the figures of the true; but into heaven itself, now to appear in the presence of God for

us: tis Iesus introiit exemplaria uerorum sed in ipsum caelum ut

9:25 Nor yet that he should offer himself often, as the high appareat nunc uultui Dei pro nobis 9:25 neque ut saepe ofpriest entereth into the holy place every year with blood of others;

ferat semet ipsum quemadmodum pontifex intrat in sancta

9:26 For then must he often have suffered since the foun- per singulos annos in sanguine alieno 9:26 alioquin oporte-

dation of the world: but now once in the end of the world hath he appeared to put away sin by the sacrifice of him-

self. bat eum frequenter pati ab origine mundi nunc autem semel

in consummatione saeculorum ad destitutionem peccati per

9:27 And as it is appointed unto men once to die, but after hostiam suam apparuit 9:27 et quemadmodum statutum est this the judgment:

Epistola Pauli Ad Hebraeos

hominibus semel mori post hoc autem iudicium 9:28 sic et 9:28 So Christ was once offered to bear the sins of many;

and unto them that look for him shall he appear the second time without sin unto salvation.

Christus semel oblatus ad multorum exhaurienda peccata secundo sine peccato apparebit expectantibus se in salutem

10:1 umbram enim habens lex bonorum futurorum non ip- 10:1 For the law having a shadow of good things to come,

and not the very image of the things, can never with those sacrifices which they offered year by year contin-

sam imaginem rerum per singulos annos hisdem ipsis hos- ually make the comers thereunto perfect. tiis quas offerunt indesinenter numquam potest accedentes perfectos facere 10:2 alioquin non cessassent offerri ideo 10:2 For then would they not have ceased to be offered?

because that the worshippers once purged should have had no more conscience of sins.

quod nullam haberent ultra conscientiam peccati cultores

semel mundati 10:3 sed in ipsis commemoratio peccato- 10:3 But in those sacrifices there is a remembrance again

made of sins every year.

rum per singulos annos fit 10:4 inpossibile enim est san- 10:4 For it is not possible that the blood of bulls and of

goats should take away sins.

guine taurorum et hircorum auferri peccata 10:5 ideo in- 10:5 Wherefore when he cometh into the world, he saith,

Sacrifice and offering thou wouldest not, but a body hast thou prepared me:

grediens mundum dicit hostiam et oblationem noluisti corpus autem aptasti mihi 10:6 holocaustomata et pro peccato 10:6 In burnt offerings and sacrifices for sin thou hast had

no pleasure.

non tibi placuit 10:7 tunc dixi ecce uenio in capitulo libri 10:7 Then said I, Lo, I come (in the volume of the book it

is written of me,) to do thy will, O God.

scriptum est de me ut faciam Deus uoluntatem tuam 10:8 10:8 Above when he said, Sacrifice and offering and burnt offerings and offering for sin thou wouldest not, neither hadst pleasure therein; which are offered by the law;

superius dicens quia hostias et oblationes et holocaustomata et pro peccato noluisti nec placita sunt tibi quae secundum legem offeruntur 10:9 tunc dixit ecce uenio ut faciam De- 10:9 Then said he, Lo, I come to do thy will, O God. He

taketh away the first, that he may establish the second.

us uoluntatem tuam aufert primum ut sequens statuat 10:10 10:10 By the which will we are sanctified through the of-

fering of the body of Jesus Christ once for all.

in qua uoluntate sanctificati sumus per oblationem corpo-

ris Christi Iesu in semel 10:11 et omnis quidem sacerdos 10:11 And every priest standeth daily ministering and offering oftentimes the same sacrifices, which can never take away sins:

praesto est cotidie ministrans et easdem saepe offerens hostias quae numquam possunt auferre peccata 10:12 hic au- 10:12 But this man, after he had offered one sacrifice for

sins for ever, sat down on the right hand of God;

tem unam pro peccatis offerens hostiam in sempiternum

sedit in dextera Dei 10:13 de cetero expectans donec po- 10:13 From henceforth expecting till his enemies be made

his footstool.

nantur inimici eius scabillum pedum eius 10:14 una enim 10:14 For by one offering he hath perfected for ever them

that are sanctified.

oblatione consummauit in sempiternum sanctificatos 10:15 10:15 Whereof the Holy Ghost also is a witness to us: for

after that he had said before,

contestatur autem nos et Spiritus Sanctus postquam enim dixit 10:16 hoc autem testamentum quod testabor ad illos 10:16 This is the covenant that I will make with them after

those days, saith the Lord, I will put my laws into their hearts, and in their minds will I write them;

post dies illos dicit Dominus dando leges meas in cordi-

bus eorum et in mente eorum superscribam eas 10:17 et 10:17 And their sins and iniquities will I remember no

more.

peccatorum et iniquitatium eorum iam non recordabor amplius 10:18 ubi autem horum remissio iam non oblatio pro 10:18 Now where remission of these is, there is no more

offering for sin.

Biblia Sacra Vulgata

10:19 Having therefore, brethren, boldness to enter into peccato 10:19 habentes itaque fratres fiduciam in introitu

the holiest by the blood of Jesus,

10:20 By a new and living way, which he hath consecrated sanctorum in sanguine Christi 10:20 quam initiauit nobis

for us, through the veil, that is to say, his flesh;

uiam nouam et uiuentem per uelamen id est carnem suam

10:21 And having an high priest over the house of God; 10:21 et sacerdotem magnum super domum Dei 10:22 acce-

10:22 Let us draw near with a true heart in full assurance of faith, having our hearts sprinkled from an evil conscience, and our bodies washed with pure water.

damus cum uero corde in plenitudine fidei aspersi corda a

10:23 Let us hold fast the profession of our faith without conscientia mala et abluti corpus aqua munda 10:23 tenea-

wavering; (for he is faithful that promised;)

mus spei nostrae confessionem indeclinabilem fidelis enim

10:24 And let us consider one another to provoke unto love est qui repromisit 10:24 et consideremus inuicem in prouo-

and to good works: 10:25 Not forsaking the assembling of ourselves together, cationem caritatis et bonorum operum 10:25 non deserentes as the manner of some is; but exhorting one another: and so much the more, as ye see the day approaching.

collectionem nostram sicut est consuetudinis quibusdam sed consolantes et tanto magis quanto uideritis adpropinquan-

10:26 For if we sin wilfully after that we have received the tem diem 10:26 uoluntarie enim peccantibus nobis post acknowledge of the truth, there remaineth no more sacrifice for sins,

ceptam notitiam ueritatis iam non relinquitur pro peccatis

10:27 But a certain fearful looking for of judgment and hostia 10:27 terribilis autem quaedam expectatio iudicii et

fiery indignation, which shall devour the adversaries.

10:28 He that despised Moses’ law died without mercy ignis aemulatio quae consumptura est aduersarios 10:28 ir-

under two or three witnesses:

ritam quis faciens legem Mosi sine ulla miseratione duobus

10:29 Of how much sorer punishment, suppose ye, shall uel tribus testibus moritur 10:29 quanto magis putatis dete-

he be thought worthy, who hath trodden under foot the Son of God, and hath counted the blood of the covenant, wherewith he was sanctified, an unholy thing, and hath done despite unto the Spirit of grace?

riora mereri supplicia qui Filium Dei conculcauerit et sanguinem testamenti pollutum duxerit in quo sanctificatus est

10:30 For we know him that hath said, Vengeance be- et Spiritui gratiae contumeliam fecerit 10:30 scimus enim

longeth unto me, I will recompense, saith the Lord. And again, The Lord shall judge his people.

qui dixit mihi uindictam ego reddam et iterum quia iudica-

10:31 It is a fearful thing to fall into the hands of the living bit Dominus populum suum 10:31 horrendum est incidere God.

10:32 But call to remembrance the former days, in which, in manus Dei uiuentis 10:32 rememoramini autem pristinos

after ye were illuminated, ye endured a great fight of afflictions;

dies in quibus inluminati magnum certamen sustinuistis pas-

10:33 Partly, whilst ye were made a gazingstock both by sionum 10:33 et in altero quidem obprobriis et tribulationi-

reproaches and afflictions; and partly, whilst ye became companions of them that were so used.

bus spectaculum facti in altero autem socii taliter conuer-

10:34 For ye had compassion of me in my bonds, and took santium effecti 10:34 nam et uinctis conpassi estis et rapi-

joyfully the spoiling of your goods, knowing in yourselves that ye have in heaven a better and an enduring substance.

nam bonorum uestrorum cum gaudio suscepistis cognoscen-

10:35 Cast not away therefore your confidence, which tes uos habere meliorem et manentem substantiam 10:35

hath great recompence of reward.

nolite itaque amittere confidentiam uestram quae magnam

10:36 For ye have need of patience, that, after ye have habet remunerationem 10:36 patientia enim uobis necessa-

done the will of God, ye might receive the promise.

ria est ut uoluntatem Dei facientes reportetis promissionem

10:37 For yet a little while, and he that shall come will 10:37 adhuc enim modicum quantulum qui uenturus est ue-

come, and will not tarry.

Epistola Pauli Ad Hebraeos

niet et non tardabit 10:38 iustus autem meus ex fide uiuit 10:38 Now the just shall live by faith: but if any man draw

back, my soul shall have no pleasure in him.

quod si subtraxerit se non placebit animae meae 10:39 nos 10:39 But we are not of them who draw back unto perdi-

tion; but of them that believe to the saving of the soul.

autem non sumus subtractionis in perditionem sed fidei in adquisitionem animae

11:1 est autem fides sperandorum substantia rerum argu- 11:1 Now faith is the substance of things hoped for, the

evidence of things not seen.

mentum non parentum 11:2 in hac enim testimonium con- 11:2 For by it the elders obtained a good report. secuti sunt senes 11:3 fide intellegimus aptata esse saecula 11:3 Through faith we understand that the worlds were

framed by the word of God, so that things which are seen were not made of things which do appear.

uerbo Dei ut ex inuisibilibus uisibilia fierent 11:4 fide plu-

11:4 By faith Abel offered unto God a more excellent sacrifice than Cain, by which he obtained witness that he was righteous, God testifying of his gifts: and by it he being dead yet speaketh.

rimam hostiam Abel quam Cain obtulit Deo per quam testimonium consecutus est esse iustus testimonium perhibente muneribus eius Deo et per illam defunctus adhuc loquitur 11:5 fide Enoch translatus est ne uideret mortem et non inue- 11:5 By faith Enoch was translated that he should not see

death; and was not found, because God had translated him: for before his translation he had this testimony, that he

niebatur quia transtulit illum Deus ante translationem enim pleased God. testimonium habebat placuisse Deo 11:6 sine fide autem in- 11:6 But without faith it is impossible to please him: for

he that cometh to God must believe that he is, and that he is a rewarder of them that diligently seek him.

possibile placere credere enim oportet accedentem ad Deum quia est et inquirentibus se remunerator fit 11:7 fide 11:7 By faith Noah, being warned of God of things not

seen as yet, moved with fear, prepared an ark to the saving of his house; by the which he condemned the world, and

Noe responso accepto de his quae adhuc non uidebantur me- became heir of the righteousness which is by faith. tuens aptauit arcam in salutem domus suae per quam damnauit mundum et iustitiae quae per fidem est heres est institutus 11:8 fide qui uocatur Abraham oboediuit in locum 11:8 By faith Abraham, when he was called to go out into

a place which he should after receive for an inheritance, obeyed; and he went out, not knowing whither he went.

exire quem accepturus erat in hereditatem et exiit nesciens quo iret 11:9 fide moratus est in terra repromissionis tam- 11:9 By faith he sojourned in the land of promise, as in

a strange country, dwelling in tabernacles with Isaac and Jacob, the heirs with him of the same promise:

quam in aliena in casulis habitando cum Isaac et Iacob coheredibus repromissionis eiusdem 11:10 expectabat enim fun- 11:10 For he looked for a city which hath foundations,

whose builder and maker is God.

damenta habentem ciuitatem cuius artifex et conditor Deus 11:11 fide et ipsa Sarra sterilis uirtutem in conceptionem se- 11:11 Through faith also Sara herself received strength

to conceive seed, and was delivered of a child when she was past age, because she judged him faithful who had

minis accepit etiam praeter tempus aetatis quoniam fidelem promised. credidit esse qui promiserat 11:12 propter quod et ab uno 11:12 Therefore sprang there even of one, and him as good

as dead, so many as the stars of the sky in multitude, and as the sand which is by the sea shore innumerable.

orti sunt et haec emortuo tamquam sidera caeli in multitudinem et sicut harena quae est ad oram maris innumerabilis 11:13 iuxta fidem defuncti sunt omnes isti non acceptis 11:13 These all died in faith, not having received the

promises, but having seen them afar off, and were persuaded of them, and embraced them, and confessed that

repromissionibus sed a longe eas aspicientes et salutantes they were strangers and pilgrims on the earth. et confitentes quia peregrini et hospites sunt supra terram

Biblia Sacra Vulgata

11:14 For they that say such things declare plainly that 11:14 qui enim haec dicunt significant se patriam inquirere

they seek a country.

11:15 And truly, if they had been mindful of that country 11:15 et si quidem illius meminissent de qua exierunt habe-

from whence they came out, they might have had opportunity to have returned.

bant utique tempus reuertendi 11:16 nunc autem meliorem

11:16 But now they desire a better country, that is, an heavenly: wherefore God is not ashamed to be called their God: for he hath prepared for them a city.

appetunt id est caelestem ideo non confunditur Deus uocari

11:17 By faith Abraham, when he was tried, offered up Deus eorum parauit enim illis ciuitatem 11:17 fide obtulit

Isaac: and he that had received the promises offered up his only begotten son,

Abraham Isaac cum temptaretur et unigenitum offerebat qui

11:18 Of whom it was said, That in Isaac shall thy seed be susceperat repromissiones 11:18 ad quem dictum est quia

called:

11:19 Accounting that God was able to raise him up, even in Isaac uocabitur tibi semen 11:19 arbitrans quia et a mor-

from the dead; from whence also he received him in a figure.

tuis suscitare potens est Deus unde eum et in parabola ac-

11:20 By faith Isaac blessed Jacob and Esau concerning cepit 11:20 fide et de futuris benedixit Isaac Iacob et Esau

things to come.

11:21 By faith Jacob, when he was a dying, blessed both 11:21 fide Iacob moriens singulis filiorum Ioseph benedixit

the sons of Joseph; and worshipped, leaning upon the top of his staff.

et adorauit fastigium uirgae eius 11:22 fide Ioseph moriens

11:22 By faith Joseph, when he died, made mention of the departing of the children of Israel; and gave commandment concerning his bones.

de profectione filiorum Israhel memoratus est et de ossibus

11:23 By faith Moses, when he was born, was hid three suis mandauit 11:23 fide Moses natus occultatus est men-

months of his parents, because they saw he was a proper child; and they were not afraid of the king’s command-

ment. sibus tribus a parentibus suis eo quod uidissent elegantem 11:24 By faith Moses, when he was come to years, refused infantem et non timuerunt regis edictum 11:24 fide Moses

to be called the son of Pharaoh’s daughter;

11:25 Choosing rather to suffer affliction with the people grandis factus negauit se esse filium filiae Pharaonis 11:25

of God, than to enjoy the pleasures of sin for a season;

magis eligens adfligi cum populo Dei quam temporalis pec-

11:26 Esteeming the reproach of Christ greater riches than cati habere iucunditatem 11:26 maiores diuitias aestimans

the treasures in Egypt: for he had respect unto the recompence of the reward.

thesauro Aegyptiorum inproperium Christi aspiciebat enim

11:27 By faith he forsook Egypt, not fearing the wrath of in remunerationem 11:27 fide reliquit Aegyptum non ueri-

the king: for he endured, as seeing him who is invisible.

tus animositatem regis inuisibilem enim tamquam uidens su-

11:28 Through faith he kept the passover, and the sprin- stinuit 11:28 fide celebrauit pascha et sanguinis effusionem

kling of blood, lest he that destroyed the firstborn should touch them.

ne qui uastabat primitiua tangeret eos 11:29 fide transierunt

11:29 By faith they passed through the Red sea as by dry land: which the Egyptians assaying to do were drowned.

mare Rubrum tamquam per aridam terram quod experti Ae-

11:30 By faith the walls of Jericho fell down, after they gyptii deuorati sunt 11:30 fide muri Hiericho ruerunt cir-

were compassed about seven days.

11:31 By faith the harlot Rahab perished not with them cuiti dierum septem 11:31 fide Raab meretrix non periit

that believed not, when she had received the spies with peace.

cum incredulis excipiens exploratores cum pace 11:32 et

11:32 And what shall I more say? for the time would fail me to tell of Gedeon, and of Barak, and of Samson, and of Jephthae; of David also, and Samuel, and of the prophets:

quid adhuc dicam deficiet enim me tempus enarrantem de Gedeon Barac Samson Iepthae Dauid et Samuhel et prophe-

11:33 Who through faith subdued kingdoms, wrought tis 11:33 qui per fidem deuicerunt regna operati sunt ius-

righteousness, obtained promises, stopped the mouths of lions.

titiam adepti sunt repromissiones obturauerunt ora leonum

11:34 Quenched the violence of fire, escaped the edge 11:34 extinxerunt impetum ignis effugerunt aciem gladii con-

of the sword, out of weakness were made strong, waxed valiant in fight, turned to flight the armies of the aliens.

Epistola Pauli Ad Hebraeos

ualuerunt de infirmitate fortes facti sunt in bello castra uerterunt exterorum 11:35 acceperunt mulieres de resurrectione 11:35 Women received their dead raised to life again: and

others were tortured, not accepting deliverance; that they might obtain a better resurrection:

mortuos suos alii autem distenti sunt non suscipientes redeeptionem ut meliorem inuenirent resurrectionem 11:36 alii 11:36 And others had trial of cruel mockings and scourg-

ings, yea, moreover of bonds and imprisonment:

uero ludibria et uerbera experti insuper et uincula et carceres 11:37 lapidati sunt secti sunt temptati sunt in occisione 11:37 They were stoned, they were sawn asunder, were

tempted, were slain with the sword: they wandered about in sheepskins and goatskins; being destitute, afflicted, tor-

gladii mortui sunt circumierunt in melotis in pellibus capri- mented; nis egentes angustiati adflicti 11:38 quibus dignus non erat 11:38 (Of whom the world was not worthy:) they wan-

dered in deserts, and in mountains, and in dens and caves of the earth.

mundus in solitudinibus errantes et montibus et speluncis et in cauernis terrae 11:39 et hii omnes testimonio fidei pro- 11:39 And these all, having obtained a good report through

faith, received not the promise:

bati non acceperunt repromissionem 11:40 Deo pro nobis 11:40 God having provided some better thing for us, that

they without us should not be made perfect.

melius aliquid prouidente ut ne sine nobis consummarentur

12:1 ideoque et nos tantam habentes inpositam nubem tes- 12:1 Wherefore seeing we also are compassed about with

so great a cloud of witnesses, let us lay aside every weight, and the sin which doth so easily beset us, and let us run

tium deponentes omne pondus et circumstans nos peccatum with patience the race that is set before us, per patientiam curramus propositum nobis certamen 12:2 12:2 Looking unto Jesus the author and finisher of our

faith; who for the joy that was set before him endured the cross, despising the shame, and is set down at the right

aspicientes in auctorem fidei et consummatorem Iesum qui hand of the throne of God. pro proposito sibi gaudio sustinuit crucem confusione contempta atque in dextera sedis Dei sedit 12:3 recogitate enim 12:3 For consider him that endured such contradiction of

sinners against himself, lest ye be wearied and faint in your minds.

eum qui talem sustinuit a peccatoribus aduersum semet ipsos contradictionem ut ne fatigemini animis uestris deficientes 12:4 nondum usque ad sanguinem restitistis aduersus pecca- 12:4 Ye have not yet resisted unto blood, striving against

sin.

tum repugnantes 12:5 et obliti estis consolationis quae uo- 12:5 And ye have forgotten the exhortation which

speaketh unto you as unto children, My son, despise not thou the chastening of the Lord, nor faint when thou art

bis tamquam filiis loquitur dicens fili mi noli neglegere di- rebuked of him: sciplinam Domini neque fatigeris dum ab eo argueris 12:6 12:6 For whom the Lord loveth he chasteneth, and scour-

geth every son whom he receiveth.

quem enim diligit Dominus castigat flagellat autem omnem filium quem recipit 12:7 in disciplina perseuerate tamquam 12:7 If ye endure chastening, God dealeth with you as with

sons; for what son is he whom the father chasteneth not?

filiis uobis offert Deus quis enim filius quem non corripit pater 12:8 quod si extra disciplinam estis cuius participes 12:8 But if ye be without chastisement, whereof all are

partakers, then are ye bastards, and not sons.

facti sunt omnes ergo adulteri et non filii estis 12:9 deinde 12:9 Furthermore we have had fathers of our flesh which

corrected us, and we gave them reverence: shall we not much rather be in subjection unto the Father of spirits, and

patres quidem carnis nostrae habuimus eruditores et reuere- live? bamur non multo magis obtemperabimus Patri spirituum et uiuemus 12:10 et illi quidem in tempore paucorum dierum 12:10 For they verily for a few days chastened us after

their own pleasure; but he for our profit, that we might be partakers of his holiness.

secundum uoluntatem suam erudiebant nos hic autem ad id

Biblia Sacra Vulgata

12:11 Now no chastening for the present seemeth to be quod utile est in recipiendo sanctificationem eius 12:11 om-

joyous, but grievous: nevertheless afterward it yieldeth the peaceable fruit of righteousness unto them which are ex-

ercised thereby. nis autem disciplina in praesenti quidem uidetur non esse

gaudii sed maeroris postea autem fructum pacatissimum ex-

12:12 Wherefore lift up the hands which hang down, and ercitatis per eam reddit iustitiae 12:12 propter quod remis-

the feeble knees;

12:13 And make straight paths for your feet, lest that sas manus et soluta genua erigite 12:13 et gressus rectos

which is lame be turned out of the way; but let it rather be healed.

facite pedibus uestris ut non claudicans erret magis autem

12:14 Follow peace with all men, and holiness, without sanetur 12:14 pacem sequimini cum omnibus et sanctimo-

which no man shall see the Lord: 12:15 Looking diligently lest any man fail of the grace of niam sine qua nemo uidebit Dominum 12:15 contemplantes God; lest any root of bitterness springing up trouble you, and thereby many be defiled;

ne quis desit gratiae Dei ne qua radix amaritudinis sursum

12:16 Lest there be any fornicator, or profane person, as germinans inpediat et per illam inquinentur multi 12:16 ne

Esau, who for one morsel of meat sold his birthright.

quis fornicator aut profanus ut Esau qui propter unam es-

12:17 For ye know how that afterward, when he would cam uendidit primitiua sua 12:17 scitote enim quoniam et

have inherited the blessing, he was rejected: for he found no place of repentance, though he sought it carefully with

tears. postea cupiens hereditare benedictionem reprobatus est non

enim inuenit paenitentiae locum quamquam cum lacrimis in-

12:18 For ye are not come unto the mount that might be quisisset eam 12:18 non enim accessistis ad tractabilem et

touched, and that burned with fire, nor unto blackness, and darkness, and tempest,

accensibilem ignem et turbinem et caliginem et procellam

12:19 And the sound of a trumpet, and the voice of words; 12:19 et tubae sonum et uocem uerborum quam qui audiewhich voice they that heard intreated that the word should not be spoken to them any more:

runt excusauerunt se ne eis fieret uerbum 12:20 non enim

12:20 (For they could not endure that which was commanded, And if so much as a beast touch the mountain, it shall be stoned, or thrust through with a dart:

portabant quod dicebatur et si bestia tetigerit montem la-

12:21 And so terrible was the sight, that Moses said, I ex- pidabitur 12:21 et ita terribile erat quod uidebatur Moses

ceedingly fear and quake:)

12:22 But ye are come unto mount Sion, and unto the city dixit exterritus sum et tremebundus 12:22 sed accessistis

of the living God, the heavenly Jerusalem, and to an innumerable company of angels,

ad Sion montem et ciuitatem Dei uiuentis Hierusalem cae-

12:23 To the general assembly and church of the firstborn, lestem et multorum milium angelorum frequentiae 12:23 et

which are written in heaven, and to God the Judge of all, and to the spirits of just men made perfect,

ecclesiam primitiuorum qui conscripti sunt in caelis et iudi-

12:24 And to Jesus the mediator of the new covenant, and cem omnium Deum et spiritus iustorum perfectorum 12:24

to the blood of sprinkling, that speaketh better things that that of Abel.

et testamenti noui mediatorem Iesum et sanguinis sparsio-

12:25 See that ye refuse not him that speaketh. For if they nem melius loquentem quam Abel 12:25 uidete ne recuescaped not who refused him that spake on earth, much more shall not we escape, if we turn away from him that

speaketh from heaven: setis loquentem si enim illi non effugerunt recusantes eum

qui super terram loquebatur multo magis nos qui de caelis

12:26 Whose voice then shook the earth: but now he hath loquentem nobis auertimur 12:26 cuius uox mouit terram

promised, saying, Yet once more I shake not the earth only, but also heaven.

tunc modo autem repromittit dicens adhuc semel ego mo-

12:27 And this word, Yet once more, signifieth the remov- uebo non solum terram sed et caelum 12:27 quod autem ing of those things that are shaken, as of things that are made, that those things which cannot be shaken may re-

main. adhuc semel dicit declarat mobilium translationem tamquam

Epistola Pauli Ad Hebraeos

factorum ut maneant ea quae sunt inmobilia 12:28 itaque 12:28 Wherefore we receiving a kingdom which cannot

be moved, let us have grace, whereby we may serve God acceptably with reverence and godly fear:

regnum inmobile suscipientes habemus gratiam per quam seruiamus placentes Deo cum metu et reuerentia 12:29 ete- 12:29 For our God is a consuming fire. nim Deus noster ignis consumens est

13:1 caritas fraternitatis maneat 13:2 hospitalitatem no- 13:1 Let brotherly love continue.

13:2 Be not forgetful to entertain strangers: for thereby some have entertained angels unawares.

lite obliuisci per hanc enim latuerunt quidam angelis hospitio receptis 13:3 mementote uinctorum tamquam simul 13:3 Remember them that are in bonds, as bound with

them; and them which suffer adversity, as being yourselves also in the body.

uincti et laborantium tamquam et ipsi in corpore morantes 13:4 honorabile conubium in omnibus et torus inmaculatus 13:4 Marriage is honourable in all, and the bed undefiled:

but whoremongers and adulterers God will judge.

fornicatores enim et adulteros iudicabit Deus 13:5 sint mo- 13:5 Let your conversation be without covetousness; and

be content with such things as ye have: for he hath said, I will never leave thee, nor forsake thee.

res sine auaritia contenti praesentibus ipse enim dixit non te deseram neque derelinquam 13:6 ita ut confidenter di- 13:6 So that we may boldly say, The Lord is my helper,

and I will not fear what man shall do unto me.

camus Dominus mihi adiutor non timebo quid faciat mihi homo 13:7 mementote praepositorum uestrorum qui uobis 13:7 Remember them which have the rule over you, who

have spoken unto you the word of God: whose faith follow, considering the end of their conversation.

locuti sunt uerbum Dei quorum intuentes exitum conuersationis imitamini fidem 13:8 Iesus Christus heri et hodie 13:8 Jesus Christ the same yesterday, and to day, and for

ever.

ipse et in saecula 13:9 doctrinis uariis et peregrinis nolite 13:9 Be not carried about with divers and strange doc-

trines. For it is a good thing that the heart be established with grace; not with meats, which have not profited them

abduci optimum enim est gratia stabiliri cor non escis quae that have been occupied therein. non profuerunt ambulantibus in eis 13:10 habemus altare de 13:10 We have an altar, whereof they have no right to eat

which serve the tabernacle.

quo edere non habent potestatem qui tabernaculo deseruiunt 13:11 quorum enim animalium infertur sanguis pro peccato 13:11 For the bodies of those beasts, whose blood is

brought into the sanctuary by the high priest for sin, are burned without the camp.

in sancta per pontificem horum corpora cremantur extra castra 13:12 propter quod et Iesus ut sanctificaret per suum 13:12 Wherefore Jesus also, that he might sanctify the

people with his own blood, suffered without the gate.

sanguinem populum extra portam passus est 13:13 exea- 13:13 Let us go forth therefore unto him without the camp,

bearing his reproach.

mus igitur ad eum extra castra inproperium eius portantes 13:14 non enim habemus hic manentem ciuitatem sed futu- 13:14 For here have we no continuing city, but we seek

one to come.

ram inquirimus 13:15 per ipsum ergo offeramus hostiam 13:15 By him therefore let us offer the sacrifice of praise to God continually, that is, the fruit of our lips giving thanks to his name.

laudis semper Deo id est fructum labiorum confitentium nomini eius 13:16 beneficientiae autem et communionis nolite 13:16 But to do good and to communicate forget not: for

with such sacrifices God is well pleased.

obliuisci talibus enim hostiis promeretur Deus 13:17 oboe- 13:17 Obey them that have the rule over you, and submit

yourselves: for they watch for your souls, as they that must give account, that they may do it with joy, and not with

dite praepositis uestris et subiacete eis ipsi enim peruigilant grief: for that is unprofitable for you. quasi rationem pro animabus uestris reddituri ut cum gaudio hoc faciant et non gementes hoc enim non expedit uobis

Biblia Sacra Vulgata

13:18 Pray for us: for we trust we have a good conscience, 13:18 orate pro nobis confidimus enim quia bonam conscien-

in all things willing to live honestly.

13:19 But I beseech you the rather to do this, that I may tiam habemus in omnibus bene uolentes conuersari 13:19

be restored to you the sooner.

amplius autem deprecor uos hoc facere ut quo celerius re-

13:20 Now the God of peace, that brought again from stituar uobis 13:20 Deus autem pacis qui eduxit de mor-

the dead our Lord Jesus, that great shepherd of the sheep, through the blood of the everlasting covenant,

tuis pastorem magnum ouium in sanguine testamenti aeterni

13:21 Make you perfect in every good work to do his will, Dominum nostrum Iesum 13:21 aptet uos in omni bono ut working in you that which is wellpleasing in his sight, through Jesus Christ; to whom be glory for ever and ever.

Amen. faciatis uoluntatem eius faciens in uobis quod placeat co-

ram se per Iesum Christum cui gloria in saecula saeculorum

13:22 And I beseech you, brethren, suffer the word of ex- amen 13:22 rogo autem uos fratres sufferatis uerbum sola-

hortation: for I have written a letter unto you in few words.

13:23 Know ye that our brother Timothy is set at liberty; cii etenim perpaucis scripsi uobis 13:23 cognoscite fratrem

with whom, if he come shortly, I will see you.

nostrum Timotheum dimissum cum quo si celerius uenerit

13:24 Salute all them that have the rule over you, and all uidebo uos 13:24 salutate omnes praepositos uestros et om-

the saints. They of Italy salute you.

13:25 Grace be with you all. Amen. nes sanctos salutant uos de Italia 13:25 gratia cum omnibus uobis amen

Epistola Catholica Iacobi

1:1 Iacobus Dei et Domini nostri Iesu Christi seruus duo- 1:1 James, a servant of God and of the Lord Jesus Christ,

to the twelve tribes which are scattered abroad, greeting.

decim tribubus quae sunt in dispersione salutem 1:2 omne 1:2 My brethren, count it all joy when ye fall into divers

temptations;

gaudium existimate fratres mei cum in temptationibus uariis incideritis 1:3 scientes quod probatio fidei uestrae patien- 1:3 Knowing this, that the trying of your faith worketh

patience.

tiam operatur 1:4 patientia autem opus perfectum habeat ut 1:4 But let patience have her perfect work, that ye may be

perfect and entire, wanting nothing.

sitis perfecti et integri in nullo deficientes 1:5 si quis autem 1:5 If any of you lack wisdom, let him ask of God, that

giveth to all men liberally, and upbraideth not; and it shall be given him.

uestrum indiget sapientiam postulet a Deo qui dat omnibus affluenter et non inproperat et dabitur ei 1:6 postulet au- 1:6 But let him ask in faith, nothing wavering. For he that

wavereth is like a wave of the sea driven with the wind and tossed.

tem in fide nihil haesitans qui enim haesitat similis est fluctui maris qui a uento mouetur et circumfertur 1:7 non ergo 1:7 For let not that man think that he shall receive any

thing of the Lord.

aestimet homo ille quod accipiat aliquid a Domino 1:8 uir 1:8 A double minded man is unstable in all his ways. duplex animo inconstans in omnibus uiis suis 1:9 glorietur 1:9 Let the brother of low degree rejoice in that he is ex-

alted:

autem frater humilis in exaltatione sua 1:10 diues autem in 1:10 But the rich, in that he is made low: because as the

flower of the grass he shall pass away.

humilitate sua quoniam sicut flos faeni transibit 1:11 exor- 1:11 For the sun is no sooner risen with a burning heat,

but it withereth the grass, and the flower thereof falleth, and the grace of the fashion of it perisheth: so also shall

tus est enim sol cum ardore et arefecit faenum et flos eius the rich man fade away in his ways. decidit et decor uultus eius deperiit ita et diues in itineribus suis marcescet 1:12 beatus uir qui suffert temptationem 1:12 Blessed is the man that endureth temptation: for

when he is tried, he shall receive the crown of life, which the Lord hath promised to them that love him.

quia cum probatus fuerit accipiet coronam uitae quam repromisit Deus diligentibus se 1:13 nemo cum temptatur dicat 1:13 Let no man say when he is tempted, I am tempted

of God: for God cannot be tempted with evil, neither tempteth he any man:

quoniam a Deo temptor Deus enim intemptator malorum est ipse autem neminem temptat 1:14 unusquisque uero tem- 1:14 But every man is tempted, when he is drawn away of

his own lust, and enticed.

ptatur a concupiscentia sua abstractus et inlectus 1:15 de- 1:15 Then when lust hath conceived, it bringeth forth sin:

and sin, when it is finished, bringeth forth death.

in concupiscentia cum conceperit parit peccatum peccatum uero cum consummatum fuerit generat mortem 1:16 nolite 1:16 Do not err, my beloved brethren. itaque errare fratres mei dilectissimi 1:17 omne datum opti- 1:17 Every good gift and every perfect gift is from above,

and cometh down from the Father of lights, with whom is no variableness, neither shadow of turning.

mum et omne donum perfectum desursum est descendens a Patre luminum apud quem non est transmutatio nec uicissitudinis obumbratio 1:18 uoluntarie genuit nos uerbo uerita- 1:18 Of his own will begat he us with the word of truth,

that we should be a kind of firstfruits of his creatures.

tis ut simus initium aliquod creaturae eius 1:19 scitis fratres 1:19 Wherefore, my beloved brethren, let every man be

swift to hear, slow to speak, slow to wrath:

mei dilecti sit autem omnis homo uelox ad audiendum tardus autem ad loquendum et tardus ad iram 1:20 ira enim uiri iu- 1:20 For the wrath of man worketh not the righteousness

of God.

stitiam Dei non operatur 1:21 propter quod abicientes om- 1:21 Wherefore lay apart all filthiness and superfluity

of naughtiness, and receive with meekness the engrafted word, which is able to save your souls.

nem inmunditiam et abundantiam malitiae in mansuetudine

Biblia Sacra Vulgata

suscipite insitum uerbum quod potest saluare animas uestras

1:22 But be ye doers of the word, and not hearers only, 1:22 estote autem factores uerbi et non auditores tantum fal-

deceiving your own selves.

1:23 For if any be a hearer of the word, and not a doer, he lentes uosmet ipsos 1:23 quia si quis auditor est uerbi et non

is like unto a man beholding his natural face in a glass:

factor hic conparabitur uiro consideranti uultum natiuitatis

1:24 For he beholdeth himself, and goeth his way, and suae in speculo 1:24 considerauit enim se et abiit et statim

straightway forgetteth what manner of man he was.

1:25 But whoso looketh into the perfect law of liberty, and oblitus est qualis fuerit 1:25 qui autem perspexerit in lege

continueth therein, he being not a forgetful hearer, but a doer of the work, this man shall be blessed in his deed.

perfecta libertatis et permanserit non auditor obliuiosus fac-

1:26 If any man among you seem to be religious, and tus sed factor operis hic beatus in facto suo erit 1:26 si quis

bridleth not his tongue, but deceiveth his own heart, this man’s religion is vain.

autem putat se religiosum esse non refrenans linguam suam

1:27 Pure religion and undefiled before God and the Father sed seducens cor suum huius uana est religio 1:27 religio is this, To visit the fatherless and widows in their affliction, and to keep himself unspotted from the world.

munda et inmaculata apud Deum et Patrem haec est uisitare pupillos et uiduas in tribulatione eorum inmaculatum se custodire ab hoc saeculo

2:1 My brethren, have not the faith of our Lord Jesus 2:1 fratres mei nolite in personarum acceptione habere fi-

Christ, the Lord of glory, with respect of persons.

2:2 For if there come unto your assembly a man with a dem Domini nostri Iesu Christi gloriae 2:2 etenim si intro-

gold ring, in goodly apparel, and there come in also a poor man in vile raiment;

ierit in conuentu uestro uir aureum anulum habens in ueste

2:3 And ye have respect to him that weareth the gay cloth- candida introierit autem et pauper in sordido habitu 2:3 et

ing, and say unto him, Sit thou here in a good place; and say to the poor, Stand thou there, or sit here under my

footstool: intendatis in eum qui indutus est ueste praeclara et dixeritis

tu sede hic bene pauperi autem dicatis tu sta illic aut sede

2:4 Are ye not then partial in yourselves, and are become sub scabillo pedum meorum 2:4 nonne iudicatis apud uosjudges of evil thoughts?

2:5 Hearken, my beloved brethren, Hath not God chosen met ipsos et facti estis iudices cogitationum iniquarum 2:5

the poor of this world rich in faith, and heirs of the kingdom which he hath promised to them that love him?

audite fratres mei dilectissimi nonne Deus elegit pauperes in hoc mundo diuites in fide et heredes regni quod repromisit

2:6 But ye have despised the poor. Do not rich men op- Deus diligentibus se 2:6 uos autem exhonorastis pauperem

press you, and draw you before the judgment seats?

nonne diuites per potentiam opprimunt uos et ipsi trahunt

2:7 Do not they blaspheme that worthy name by the which uos ad iudicia 2:7 nonne ipsi blasphemant bonum nomen ye are called?

2:8 If ye fulfil the royal law according to the scripture, quod inuocatum est super uos 2:8 si tamen legem perficitis

Thou shalt love thy neighbour as thyself, ye do well:

regalem secundum scripturas diliges proximum tuum sicut

2:9 But if ye have respect to persons, ye commit sin, and te ipsum bene facitis 2:9 si autem personas accipitis pec-

are convinced of the law as transgressors.

2:10 For whosoever shall keep the whole law, and yet of- catum operamini redarguti a lege quasi transgressores 2:10

fend in one point, he is guilty of all.

quicumque autem totam legem seruauerit offendat autem in

2:11 For he that said, Do not commit adultery, said also, uno factus est omnium reus 2:11 qui enim dixit non moe-Do not kill. Now if thou commit no adultery, yet if thou kill, thou art become a transgressor of the law.

chaberis dixit et non occides quod si non moechaberis occi-

Epistola Catholica Iacobi

des autem factus es transgressor legis 2:12 sic loquimini et 2:12 So speak ye, and so do, as they that shall be judged

by the law of liberty.

sic facite sicut per legem libertatis incipientes iudicari 2:13 2:13 For he shall have judgment without mercy, that hath

shewed no mercy; and mercy rejoiceth against judgment.

iudicium enim sine misericordia illi qui non fecit misericordiam superexultat autem misericordia iudicio 2:14 quid 2:14 What doth it profit, my brethren, though a man say

he hath faith, and have not works? can faith save him?

prodeeit fratres mei si fidem quis dicat se habere opera au-

tem non habeat numquid poterit fides saluare eum 2:15 si 2:15 If a brother or sister be naked, and destitute of daily

food,

autem frater aut soror nudi sunt et indigent uictu cotidiano 2:16 dicat autem aliquis de uobis illis ite in pace calefaci- 2:16 And one of you say unto them, Depart in peace, be

ye warmed and filled; notwithstanding ye give them not those things which are needful to the body; what doth it

mini et saturamini non dederitis autem eis quae necessaria profit? sunt corporis quid prodeeit 2:17 sic et fides si non habeat 2:17 Even so faith, if it hath not works, is dead, being

alone.

opera mortua est in semet ipsam 2:18 sed dicet quis tu fi- 2:18 Yea, a man may say, Thou hast faith, and I have

works: shew me thy faith without thy works, and I will shew thee my faith by my works.

dem habes et ego opera habeo ostende mihi fidem tuam sine

operibus et ego ostendam tibi ex operibus fidem meam 2:19 2:19 Thou believest that there is one God; thou doest well:

the devils also believe, and tremble.

tu credis quoniam unus est Deus bene facis et daemones credunt et contremescunt 2:20 uis autem scire o homo inanis 2:20 But wilt thou know, O vain man, that faith without

works is dead?

quoniam fides sine operibus otiosa est 2:21 Abraham pa- 2:21 Was not Abraham our father justified by works, when

he had offered Isaac his son upon the altar?

ter noster nonne ex operibus iustificatus est offerens Isaac filium suum super altare 2:22 uides quoniam fides coope- 2:22 Seest thou how faith wrought with his works, and by

works was faith made perfect?

rabatur operibus illius et ex operibus fides consummata est 2:23 et suppleta est scriptura dicens credidit Abraham Deo et 2:23 And the scripture was fulfilled which saith, Abraham

believed God, and it was imputed unto him for righteousness: and he was called the Friend of God.

reputatum est illi ad iustitiam et amicus Dei appellatus est 2:24 uidetis quoniam ex operibus iustificatur homo et non ex 2:24 Ye see then how that by works a man is justified, and

not by faith only.

fide tantum 2:25 similiter autem et Raab meretrix nonne ex 2:25 Likewise also was not Rahab the harlot justified by

works, when she had received the messengers, and had sent them out another way?

operibus iustificata est suscipiens nuntios et alia uia eiciens 2:26 sicut enim corpus sine spiritu emortuum est ita et fides 2:26 For as the body without the spirit is dead, so faith

without works is dead also.

sine operibus mortua est

3:1 nolite plures magistri fieri fratres mei scientes quoniam 3:1 My brethren, be not many masters, knowing that we

shall receive the greater condemnation.

maius iudicium sumitis 3:2 in multis enim offendimus om- 3:2 For in many things we offend all. If any man offend

not in word, the same is a perfect man, and able also to bridle the whole body.

nes si quis in uerbo non offendit hic perfectus est uir potens etiam freno circumducere totum corpus 3:3 si autem equo- 3:3 Behold, we put bits in the horses’ mouths, that they

may obey us; and we turn about their whole body.

rum frenos in ora mittimus ad consentiendum nobis et omne corpus illorum circumferimus 3:4 ecce et naues cum mag- 3:4 Behold also the ships, which though they be so great,

and are driven of fierce winds, yet are they turned about with a very small helm, whithersoever the governor listeth.

nae sint et a uentis ualidis minentur circumferuntur a modico

Biblia Sacra Vulgata

3:5 Even so the tongue is a little member, and boast- gubernaculo ubi impetus dirigentis uoluerit 3:5 ita et lin-

eth great things. Behold, how great a matter a little fire kindleth!

gua modicum quidem membrum est et magna exultat ecce

3:6 And the tongue is a fire, a world of iniquity: so is quantus ignis quam magnam siluam incendit 3:6 et lingua

the tongue among our members, that it defileth the whole body, and setteth on fire the course of nature; and it is set

on fire of hell. ignis est uniuersitas iniquitatis lingua constituitur in mem-

bris nostris quae maculat totum corpus et inflammat rotam

3:7 For every kind of beasts, and of birds, and of serpents, natiuitatis nostrae inflammata a gehenna 3:7 omnis enim and of things in the sea, is tamed, and hath been tamed of mankind:

natura bestiarum et uolucrum et serpentium etiam ceterorum

3:8 But the tongue can no man tame; it is an unruly evil, domantur et domita sunt a natura humana 3:8 linguam au-

full of deadly poison.

tem nullus hominum domare potest inquietum malum plena

3:9 Therewith bless we God, even the Father; and there- ueneno mortifero 3:9 in ipsa benedicimus Dominum et Pa-

with curse we men, which are made after the similitude of God.

trem et in ipsa maledicimus homines qui ad similitudinem

3:10 Out of the same mouth proceedeth blessing and curs- Dei facti sunt 3:10 ex ipso ore procedit benedictio et ma-

ing. My brethren, these things ought not so to be.

3:11 Doth a fountain send forth at the same place sweet ledictio non oportet fratres mei haec ita fieri 3:11 numquid

water and bitter?

fons de eodem foramine emanat dulcem et amaram aquam

3:12 Can the fig tree, my brethren, bear olive berries? ei- 3:12 numquid potest fratres mei ficus oliuas facere aut uitis

ther a vine, figs? so can no fountain both yield salt water and fresh.

ficus sic neque salsa dulcem potest facere aquam 3:13 quis

3:13 Who is a wise man and endued with knowledge among you? let him shew out of a good conversation his works with meekness of wisdom.

sapiens et disciplinatus inter uos ostendat ex bona conuer-

3:14 But if ye have bitter envying and strife in your hearts, satione operationem suam in mansuetudine sapientiae 3:14 glory not, and lie not against the truth.

quod si zelum amarum habetis et contentiones in cordibus uestris nolite gloriari et mendaces esse aduersus ueritatem

3:15 This wisdom descendeth not from above, but is 3:15 non est ista sapientia desursum descendens sed terrena

earthly, sensual, devilish.

3:16 For where envying and strife is, there is confusion animalis diabolica 3:16 ubi enim zelus et contentio ibi in-

and every evil work.

3:17 But the wisdom that is from above is first pure, then constantia et omne opus prauum 3:17 quae autem desursum

peaceable, gentle, and easy to be intreated, full of mercy and good fruits, without partiality, and without hypocrisy.

est sapientia primum quidem pudica est deinde pacifica modesta suadibilis plena misericordia et fructibus bonis non iu-

3:18 And the fruit of righteousness is sown in peace of dicans sine simulatione 3:18 fructus autem iustitiae in pace

them that make peace.

seminatur facientibus pacem

4:1 From whence come wars and fightings among you? 4:1 unde bella et lites in uobis nonne hinc ex concupiscen-

come they not hence, even of your lusts that war in your members?

tiis uestris quae militant in membris uestris 4:2 concupisci-

4:2 Ye lust, and have not: ye kill, and desire to have, and cannot obtain: ye fight and war, yet ye have not, because ye ask not.

tis et non habetis occiditis et zelatis et non potestis adipisci litigatis et belligeratis non habetis propter quod non post-

4:3 Ye ask, and receive not, because ye ask amiss, that ye ulatis 4:3 petitis et non accipitis eo quod male petatis ut in

may consume it upon your lusts.

4:4 Ye adulterers and adulteresses, know ye not that the concupiscentiis uestris insumatis 4:4 adulteri nescitis quia

friendship of the world is enmity with God? whosoever therefore will be a friend of the world is the enemy of God.

Epistola Catholica Iacobi

amicitia huius mundi inimica est Dei quicumque ergo uoluerit amicus esse saeculi huius inimicus Dei constituitur 4:5 4:5 Do ye think that the scripture saith in vain, The spirit

that dwelleth in us lusteth to envy?

aut putatis quia inaniter scriptura dicat ad inuidiam concupiscit Spiritus qui inhabitat in nobis 4:6 maiorem autem dat 4:6 But he giveth more grace. Wherefore he saith, God

resisteth the proud, but giveth grace unto the humble.

gratiam propter quod dicit Deus superbis resistit humilibus autem dat gratiam 4:7 subditi igitur estote Deo resistite au- 4:7 Submit yourselves therefore to God. Resist the devil,

and he will flee from you.

tem diabolo et fugiet a uobis 4:8 adpropiate Domino et ad- 4:8 Draw nigh to God, and he will draw nigh to you.

Cleanse your hands, ye sinners; and purify your hearts, ye double minded.

propinquabit uobis emundate manus peccatores et purificate corda duplices animo 4:9 miseri estote et lugete et plorate 4:9 Be afflicted, and mourn, and weep: let your laughter

be turned to mourning, and your joy to heaviness.

risus uester in luctum conuertatur et gaudium in maerorem 4:10 humiliamini in conspectu Domini et exaltabit uos 4:11 4:10 Humble yourselves in the sight of the Lord, and he

shall lift you up.

4:11 Speak not evil one of another, brethren. He that speaketh evil of his brother, and judgeth his brother, speaketh evil of the law, and judgeth the law: but if thou judge the law, thou art not a doer of the law, but a judge.

nolite detrahere de alterutrum fratres qui detrahit fratri aut qui iudicat fratrem suum detrahit legi et iudicat legem si autem iudicas legem non es factor legis sed iudex 4:12 unus 4:12 There is one lawgiver, who is able to save and to de-

stroy: who art thou that judgest another?

est legislator et iudex qui potest perdere et liberare tu autem quis es qui iudicas proximum 4:13 ecce nunc qui dicitis ho- 4:13 Go to now, ye that say, To day or to morrow we will

go into such a city, and continue there a year, and buy and sell, and get gain:

die aut crastino ibimus in illam ciuitatem et faciemus quidem ibi annum et mercabimur et lucrum faciemus 4:14 qui igno- 4:14 Whereas ye know not what shall be on the morrow.

For what is your life? It is even a vapour, that appeareth for a little time, and then vanisheth away.

ratis quid erit in crastinum quae enim est uita uestra uapor est ad modicum parens deinceps exterminatur 4:15 pro eo 4:15 For that ye ought to say, If the Lord will, we shall

live, and do this, or that.

ut dicatis si Dominus uoluerit et uixerimus faciemus hoc aut illud 4:16 nunc autem exultatis in superbiis uestris omnis 4:16 But now ye rejoice in your boastings: all such rejoic-

ing is evil.

exultatio talis maligna est 4:17 scienti igitur bonum facere 4:17 Therefore to him that knoweth to do good, and doeth

it not, to him it is sin.

et non facienti peccatum est illi

5:1 age nunc diuites plorate ululantes in miseriis quae ad- 5:1 Go to now, ye rich men, weep and howl for your mis-

eries that shall come upon you.

uenient uobis 5:2 diuitiae uestrae putrefactae sunt et ues- 5:2 Your riches are corrupted, and your garments are

motheaten.

timenta uestra a tineis comesta sunt 5:3 aurum et argen- 5:3 Your gold and silver is cankered; and the rust of them

shall be a witness against you, and shall eat your flesh as it were fire. Ye have heaped treasure together for the last

tum uestrum eruginauit et erugo eorum in testimonium uobis days. erit et manducabit carnes uestras sicut ignis thesaurizastis in nouissimis diebus 5:4 ecce merces operariorum qui mes- 5:4 Behold, the hire of the labourers who have reaped

down your fields, which is of you kept back by fraud, crieth: and the cries of them which have reaped are entered

suerunt regiones uestras qui fraudatus est a uobis clamat et into the ears of the Lord of sabaoth. clamor ipsorum in aures Domini Sabaoth introiit 5:5 epu- 5:5 Ye have lived in pleasure on the earth, and been wan-

ton; ye have nourished your hearts, as in a day of slaughter.

lati estis super terram et in luxuriis enutristis corda uestra in

Biblia Sacra Vulgata

5:6 Ye have condemned and killed the just; and he doth die occisionis 5:6 addixistis occidistis iustum non resistit

not resist you.

5:7 Be patient therefore, brethren, unto the coming of the uobis 5:7 patientes igitur estote fratres usque ad aduentum

Lord. Behold, the husbandman waiteth for the precious fruit of the earth, and hath long patience for it, until he

receive the early and latter rain. Domini ecce agricola expectat pretiosum fructum terrae pa-

5:8 Be ye also patient; stablish your hearts: for the coming tienter ferens donec accipiat temporiuum et serotinum 5:8 of the Lord draweth nigh.

patientes estote et uos confirmate corda uestra quoniam ad-

5:9 Grudge not one against another, brethren, lest ye be uentus Domini adpropinquauit 5:9 nolite ingemescere fra-

condemned: behold, the judge standeth before the door.

tres in alterutrum ut non iudicemini ecce iudex ante ianuam

5:10 Take, my brethren, the prophets, who have spoken in adsistit 5:10 exemplum accipite fratres laboris et patientiae

the name of the Lord, for an example of suffering affliction, and of patience.

prophetas qui locuti sunt in nomine Domini 5:11 ecce bea-

5:11 Behold, we count them happy which endure. Ye have heard of the patience of Job, and have seen the end of the Lord; that the Lord is very pitiful, and of tender mercy.

tificamus qui sustinuerunt sufferentiam Iob audistis et finem Domini uidistis quoniam misericors est Dominus et misera-

5:12 But above all things, my brethren, swear not, neither tor 5:12 ante omnia autem fratres mei nolite iurare neque by heaven, neither by the earth, neither by any other oath: but let your yea be yea; and your nay, nay; lest ye fall into

condemnation. per caelum neque per terram neque aliud quodcumque iura-

mentum sit autem uestrum est est non non uti non sub iudicio

5:13 Is any among you afflicted? let him pray. Is any decidatis 5:13 tristatur aliquis uestrum oret aequo animo est

merry? let him sing psalms.

5:14 Is any sick among you? let him call for the elders psallat 5:14 infirmatur quis in uobis inducat presbyteros ec-

of the church; and let them pray over him, anointing him with oil in the name of the Lord:

clesiae et orent super eum unguentes eum oleo in nomine

5:15 And the prayer of faith shall save the sick, and the Domini 5:15 et oratio fidei saluabit infirmum et adleuabit

Lord shall raise him up; and if he have committed sins, they shall be forgiven him.

eum Dominus et si in peccatis sit dimittentur ei 5:16 confi-

5:16 Confess your faults one to another, and pray one for another, that ye may be healed. The effectual fervent prayer of a righteous man availeth much.

temini ergo alterutrum peccata uestra et orate pro inuicem ut

5:17 Elias was a man subject to like passions as we are, saluemini multum enim ualet deprecatio iusti adsidua 5:17

and he prayed earnestly that it might not rain: and it rained not on the earth by the space of three years and six months.

Helias homo erat similis nobis passibilis et oratione orauit ut non plueret super terram et non pluit annos tres et men-

5:18 And he prayed again, and the heaven gave rain, and ses sex 5:18 et rursum orauit et caelum dedit pluuiam et

the earth brought forth her fruit.

5:19 Brethren, if any of you do err from the truth, and one terra dedit fructum suum 5:19 fratres mei si quis ex uobis

convert him;

5:20 Let him know, that he which converteth the sinner errauerit a ueritate et conuerterit quis eum 5:20 scire debet

from the error of his way shall save a soul from death, and shall hide a multitude of sins.

quoniam qui conuerti fecerit peccatorem ab errore uiae suae saluabit animam eius a morte et operit multitudinem peccatorum

Epistola Petri I

1:1 Petrus apostolus Iesu Christi electis aduenis dispersio- 1:1 Peter, an apostle of Jesus Christ, to the strangers scat-

tered throughout Pontus, Galatia, Cappadocia, Asia, and Bithynia,

nis Ponti Galatiae Cappadociae Asiae et Bithyniae 1:2 se-

1:2 Elect according to the foreknowledge of God the Father, through sanctification of the Spirit, unto obedience and sprinkling of the blood of Jesus Christ: Grace unto you, and peace, be multiplied.

cundum praescientiam Dei Patris in sanctificatione Spiritus in oboedientiam et aspersionem sanguinis Iesu Christi gratia uobis et pax multiplicetur 1:3 benedictus Deus et Pater 1:3 Blessed be the God and Father of our Lord Jesus

Christ, which according to his abundant mercy hath begotten us again unto a lively hope by the resurrection of

Domini nostri Iesu Christi qui secundum magnam miseri- Jesus Christ from the dead, cordiam suam regenerauit nos in spem uiuam per resurrectionem Iesu Christi ex mortuis 1:4 in hereditatem incorrup- 1:4 To an inheritance incorruptible, and undefiled, and that

fadeth not away, reserved in heaven for you,

tibilem et incontaminatam et inmarcescibilem conseruatam in caelis in uobis 1:5 qui in uirtute Dei custodimini per fi- 1:5 Who are kept by the power of God through faith unto

salvation ready to be revealed in the last time.

dem in salutem paratam reuelari in tempore nouissimo 1:6 1:6 Wherein ye greatly rejoice, though now for a season, if

need be, ye are in heaviness through manifold temptations:

in quo exultatis modicum nunc si oportet contristati in uariis temptationibus 1:7 ut probatum uestrae fidei multo pretio- 1:7 That the trial of your faith, being much more precious

than of gold that perisheth, though it be tried with fire, might be found unto praise and honour and glory at the

sius sit auro quod perit per ignem probato inueniatur in lau- appearing of Jesus Christ: dem et gloriam et honorem in reuelatione Iesu Christi 1:8 1:8 Whom having not seen, ye love; in whom, though now

ye see him not, yet believing, ye rejoice with joy unspeakable and full of glory:

quem cum non uideritis diligitis in quem nunc quoque non uidentes credentes autem exultatis laetitia inenarrabili et glorificata 1:9 reportantes finem fidei uestrae salutem anima- 1:9 Receiving the end of your faith, even the salvation of

your souls.

rum 1:10 de qua salute exquisierunt atque scrutati sunt pro- 1:10 Of which salvation the prophets have enquired and

searched diligently, who prophesied of the grace that should come unto you:

phetae qui de futura in uobis gratia prophetauerunt 1:11 sc-

1:11 Searching what, or what manner of time the Spirit of Christ which was in them did signify, when it testified beforehand the sufferings of Christ, and the glory that should follow.

rutantes in quod uel quale tempus significaret in eis Spiritus Christi praenuntians eas quae in Christo sunt passiones et posteriores glorias 1:12 quibus reuelatum est quia non sibi 1:12 Unto whom it was revealed, that not unto themselves,

but unto us they did minister the things, which are now reported unto you by them that have preached the gospel unto you with the Holy Ghost sent down from heaven; which things the angels desire to look into.

ipsis uobis autem ministrabant ea quae nunc nuntiata sunt uobis per eos qui euangelizauerunt uos Spiritu Sancto misso de caelo in quae desiderant angeli prospicere 1:13 propter 1:13 Wherefore gird up the loins of your mind, be sober,

and hope to the end for the grace that is to be brought unto you at the revelation of Jesus Christ;

quod succincti lumbos mentis uestrae sobrii perfecte sperate in eam quae offertur uobis gratiam in reuelatione Iesu Christi 1:14 quasi filii oboedientiae non configurati prioribus 1:14 As obedient children, not fashioning yourselves ac-

cording to the former lusts in your ignorance:

ignorantiae uestrae desideriis 1:15 sed secundum eum qui 1:15 But as he which hath called you is holy, so be ye holy

in all manner of conversation;

uocauit uos sanctum et ipsi sancti in omni conuersatione sitis 1:16 quoniam scriptum est sancti eritis quia ego sanctus sum 1:16 Because it is written, Be ye holy; for I am holy. 1:17 et si Patrem inuocatis eum qui sine acceptione persona- 1:17 And if ye call on the Father, who without respect of

persons judgeth according to every man’s work, pass the time of your sojourning here in fear:

Biblia Sacra Vulgata

rum iudicat secundum uniuscuiusque opus in timore inco-

1:18 Forasmuch as ye know that ye were not redeemed latus uestri tempore conuersamini 1:18 scientes quod non

with corruptible things, as silver and gold, from your vain conversation received by tradition from your fathers;

corruptibilibus argento uel auro redeepti estis de uana ues-

1:19 But with the precious blood of Christ, as of a lamb tra conuersatione paternae traditionis 1:19 sed pretioso san-

without blemish and without spot:

1:20 Who verily was foreordained before the foundation guine quasi agni incontaminati et inmaculati Christi 1:20

of the world, but was manifest in these last times for you,

praecogniti quidem ante constitutionem mundi manifestati

1:21 Who by him do believe in God, that raised him up autem nouissimis temporibus propter uos 1:21 qui per ip-

from the dead, and gave him glory; that your faith and hope might be in God.

sum fideles estis in Deo qui suscitauit eum a mortuis et dedit

1:22 Seeing ye have purified your souls in obeying ei gloriam ut fides uestra et spes esset in Deo 1:22 animas

the truth through the Spirit unto unfeigned love of the brethren, see that ye love one another with a pure heart

fervently: uestras castificantes in oboedientia caritatis in fraternitatis 1:23 Being born again, not of corruptible seed, but of in- amore simplici ex corde inuicem diligite adtentius 1:23 re-

corruptible, by the word of God, which liveth and abideth for ever.

nati non ex semine corruptibili sed incorruptibili per uerbum

1:24 For all flesh is as grass, and all the glory of man as Dei uiui et permanentis 1:24 quia omnis caro ut faenum et

the flower of grass. The grass withereth, and the flower thereof falleth away:

omnis gloria eius tamquam flos faeni exaruit faenum et flos

1:25 But the word of the Lord endureth for ever. And this decidit 1:25 uerbum autem Domini manet in aeternum hoc is the word which by the gospel is preached unto you.

est autem uerbum quod euangelizatum est in uos

2:1 Wherefore laying aside all malice, and all guile, and 2:1 deponentes igitur omnem malitiam et omnem dolum hypocrisies, and envies, all evil speakings,

2:2 As newborn babes, desire the sincere milk of the word, et simulationes et inuidias et omnes detractiones 2:2 sicut

that ye may grow thereby:

modo geniti infantes rationale sine dolo lac concupiscite ut

2:3 If so be ye have tasted that the Lord is gracious. in eo crescatis in salutem 2:3 si gustastis quoniam dulcis

2:4 To whom coming, as unto a living stone, disallowed Dominus 2:4 ad quem accedentes lapidem uiuum ab hoindeed of men, but chosen of God, and precious,

minibus quidem reprobatum a Deo autem electum honorifi-

2:5 Ye also, as lively stones, are built up a spiritual house, catum 2:5 et ipsi tamquam lapides uiui superaedificamini an holy priesthood, to offer up spiritual sacrifices, acceptable to God by Jesus Christ.

domus spiritalis sacerdotium sanctum offerre spiritales hos-

2:6 Wherefore also it is contained in the scripture, Behold, tias acceptabiles Deo per Iesum Christum 2:6 propter quod I lay in Sion a chief corner stone, elect, precious: and he that believeth on him shall not be confounded.

continet in scriptura ecce pono in Sion lapidem summum angularem electum pretiosum et qui crediderit in eo non con-

2:7 Unto you therefore which believe he is precious: but fundetur 2:7 uobis igitur honor credentibus non credentibus

unto them which be disobedient, the stone which the builders disallowed, the same is made the head of the cor-

ner, autem lapis quem reprobauerunt aedificantes hic factus est in 2:8 And a stone of stumbling, and a rock of offence, even caput anguli 2:8 et lapis offensionis et petra scandali qui of-

to them which stumble at the word, being disobedient: whereunto also they were appointed.

fendunt uerbo nec credunt in quod et positi sunt 2:9 uos au-

2:9 But ye are a chosen generation, a royal priesthood, an holy nation, a peculiar people; that ye should shew forth the praises of him who hath called you out of darkness into his marvellous light;

tem genus electum regale sacerdotium gens sancta populus adquisitionis ut uirtutes adnuntietis eius qui de tenebris uos

Epistola Petri I

uocauit in admirabile lumen suum 2:10 qui aliquando non 2:10 Which in time past were not a people, but are now

the people of God: which had not obtained mercy, but now have obtained mercy.

populus nunc autem populus Dei qui non consecuti misericordiam nunc autem misericordiam consecuti 2:11 caris- 2:11 Dearly beloved, I beseech you as strangers and pil-

grims, abstain from fleshly lusts, which war against the soul;

simi obsecro tamquam aduenas et peregrinos abstinere uos a carnalibus desideriis quae militant aduersus animam 2:12 2:12 Having your conversation honest among the Gentiles:

that, whereas they speak against you as evildoers, they may by your good works, which they shall behold, glo-

conuersationem uestram inter gentes habentes bonam ut in rify God in the day of visitation. eo quod detractant de uobis tamquam de malefactoribus ex bonis operibus considerantes glorificent Deum in die uisitationis 2:13 subiecti estote omni humanae creaturae pro- 2:13 Submit yourselves to every ordinance of man for the

Lord’s sake: whether it be to the king, as supreme;

pter Dominum siue regi quasi praecellenti 2:14 siue duci- 2:14 Or unto governors, as unto them that are sent by him

for the punishment of evildoers, and for the praise of them that do well.

bus tamquam ab eo missis ad uindictam malefactorum lau-

dem uero bonorum 2:15 quia sic est uoluntas Dei ut be- 2:15 For so is the will of God, that with well doing ye may

put to silence the ignorance of foolish men:

nefacientes obmutescere faciatis inprudentium hominum ignorantiam 2:16 quasi liberi et non quasi uelamen habentes 2:16 As free, and not using your liberty for a cloke of ma-

liciousness, but as the servants of God.

malitiae libertatem sed sicut serui Dei 2:17 omnes honorate 2:17 Honour all men. Love the brotherhood. Fear God.

Honour the king.

fraternitatem diligite Deum timete regem honorificate 2:18 2:18 Servants, be subject to your masters with all fear; not

only to the good and gentle, but also to the froward.

serui subditi in omni timore dominis non tantum bonis et modestis sed etiam discolis 2:19 haec est enim gratia si pro- 2:19 For this is thankworthy, if a man for conscience to-

ward God endure grief, suffering wrongfully.

pter conscientiam Dei sustinet quis tristitias patiens iniuste 2:20 quae enim gloria est si peccantes et colaphizati suffer- 2:20 For what glory is it, if, when ye be buffeted for your

faults, ye shall take it patiently? but if, when ye do well, and suffer for it, ye take it patiently, this is acceptable with

tis sed si benefacientes et patientes sustinetis haec est gratia God.

apud Deum 2:21 in hoc enim uocati estis quia et Christus 2:21 For even hereunto were ye called: because Christ also suffered for us, leaving us an example, that ye should follow his steps:

passus est pro uobis uobis relinquens exemplum ut sequamini uestigia eius 2:22 qui peccatum non fecit nec inuentus 2:22 Who did no sin, neither was guile found in his mouth: est dolus in ore ipsius 2:23 qui cum malediceretur non ma- 2:23 Who, when he was reviled, reviled not again; when

he suffered, he threatened not; but committed himself to him that judgeth righteously:

ledicebat cum pateretur non comminabatur tradebat autem iudicanti se iniuste 2:24 qui peccata nostra ipse pertulit in 2:24 Who his own self bare our sins in his own body on

the tree, that we, being dead to sins, should live unto righteousness: by whose stripes ye were healed.

corpore suo super lignum ut peccatis mortui iustitiae uiueremus cuius liuore sanati estis 2:25 eratis enim sicut oues 2:25 For ye were as sheep going astray; but are now re-

turned unto the Shepherd and Bishop of your souls.

errantes sed conuersi estis nunc ad pastorem et episcopum animarum uestrarum

3:1 similiter mulieres subditae suis uiris ut et si qui non 3:1 Likewise, ye wives, be in subjection to your own hus-

bands; that, if any obey not the word, they also may without the word be won by the conversation of the wives;

credunt uerbo per mulierum conuersationem sine uerbo lucri

Biblia Sacra Vulgata

3:2 While they behold your chaste conversation coupled fiant 3:2 considerantes in timore castam conuersationem

with fear.

3:3 Whose adorning let it not be that outward adorning of uestram 3:3 quarum sit non extrinsecus capillaturae aut cir-

plaiting the hair, and of wearing of gold, or of putting on of apparel;

cumdatio auri aut indumenti uestimentorum cultus 3:4 sed

3:4 But let it be the hidden man of the heart, in that which is not corruptible, even the ornament of a meek and quiet spirit, which is in the sight of God of great price.

qui absconditus cordis est homo in incorruptibilitate quieti et

3:5 For after this manner in the old time the holy women modesti spiritus quod est in conspectu Dei locuples 3:5 sic

also, who trusted in God, adorned themselves, being in subjection unto their own husbands:

enim aliquando et sanctae mulieres sperantes in Deo orna-

3:6 Even as Sara obeyed Abraham, calling him lord: bant se subiectae propriis uiris 3:6 sicut Sarra oboediebat

whose daughters ye are, as long as ye do well, and are not afraid with any amazement.

Abrahae dominum eum uocans cuius estis filiae benefacien-

3:7 Likewise, ye husbands, dwell with them according tes et non timentes ullam perturbationem 3:7 uiri similiter

to knowledge, giving honour unto the wife, as unto the weaker vessel, and as being heirs together of the grace of

life; that your prayers be not hindered. cohabitantes secundum scientiam quasi infirmiori uaso mu-

liebri inpertientes honorem tamquam et coheredibus gratiae

3:8 Finally, be ye all of one mind, having compassion one uitae uti ne inpediantur orationes uestrae 3:8 in fine autem

of another, love as brethren, be pitiful, be courteous:

omnes unianimes conpatientes fraternitatis amatores mise-

3:9 Not rendering evil for evil, or railing for railing: ricordes humiles 3:9 non reddentes malum pro malo uel but contrariwise blessing; knowing that ye are thereunto called, that ye should inherit a blessing.

maledictum pro maledicto sed e contrario benedicentes quia in hoc uocati estis ut benedictionem hereditate possideatis

3:10 For he that will love life, and see good days, let him 3:10 qui enim uult uitam diligere et uidere dies bonos coer-

refrain his tongue from evil, and his lips that they speak no guile:

ceat linguam suam a malo et labia eius ne loquantur dolum

3:11 Let him eschew evil, and do good; let him seek peace, 3:11 declinet autem a malo et faciat bonum inquirat pacem et

and ensue it.

3:12 For the eyes of the Lord are over the righteous, and persequatur eam 3:12 quia oculi Domini super iustos et au-

his ears are open unto their prayers: but the face of the Lord is against them that do evil.

res eius in preces eorum uultus autem Domini super facien-

3:13 And who is he that will harm you, if ye be followers tes mala 3:13 et quis est qui uobis noceat si boni aemulato-

of that which is good?

3:14 But and if ye suffer for righteousness’ sake, happy are res fueritis 3:14 sed et si quid patimini propter iustitiam be-

ye: and be not afraid of their terror, neither be troubled;

ati timorem autem eorum ne timueritis et non conturbemini

3:15 But sanctify the Lord God in your hearts: and be 3:15 Dominum autem Christum sanctificate in cordibus ue-

ready always to give an answer to every man that asketh you a reason of the hope that is in you with meekness and

fear: stris parati semper ad satisfactionem omni poscenti uos ra-

3:16 Having a good conscience; that, whereas they speak tionem de ea quae in uobis est spe 3:16 sed cum modestia evil of you, as of evildoers, they may be ashamed that falsely accuse your good conversation in Christ.

et timore conscientiam habentes bonam ut in eo quod detrahunt uobis confundantur qui calumniantur uestram bonam in

3:17 For it is better, if the will of God be so, that ye suffer Christo conuersationem 3:17 melius est enim benefacientes

for well doing, than for evil doing.

3:18 For Christ also hath once suffered for sins, the just si uelit uoluntas Dei pati quam malefacientes 3:18 quia et

for the unjust, that he might bring us to God, being put to death in the flesh, but quickened by the Spirit:

Christus semel pro peccatis mortuus est iustus pro iniustis ut nos offerret Deo mortificatus carne uiuificatus autem spiritu

Epistola Petri I

3:19 in quo et his qui in carcere erant spiritibus ueniens prae- 3:19 By which also he went and preached unto the spirits

in prison;

dicauit 3:20 qui increduli fuerant aliquando quando expec- 3:20 Which sometime were disobedient, when once the

longsuffering of God waited in the days of Noah, while the ark was a preparing, wherein few, that is, eight souls

tabat Dei patientia in diebus Noe cum fabricaretur arca in were saved by water. qua pauci id est octo animae saluae factae sunt per aquam 3:21 quod et uos nunc similis formae saluos facit baptisma 3:21 The like figure whereunto even baptism doth also

now save us (not the putting away of the filth of the flesh, but the answer of a good conscience toward God,) by the

non carnis depositio sordium sed conscientiae bonae inter- resurrection of Jesus Christ: rogatio in Deum per resurrectionem Iesu Christi 3:22 qui 3:22 Who is gone into heaven, and is on the right hand of

God; angels and authorities and powers being made subject unto him.

est in dextera Dei profectus in caelum subiectis sibi angelis et potestatibus et uirtutibus

4:1 Christo igitur passo in carne et uos eadem cogitatione 4:1 Forasmuch then as Christ hath suffered for us in the

flesh, arm yourselves likewise with the same mind: for he that hath suffered in the flesh hath ceased from sin;

armamini quia qui passus est carne desiit a peccatis 4:2 ut

4:2 That he no longer should live the rest of his time in the flesh to the lusts of men, but to the will of God.

iam non hominum desideriis sed uoluntate Dei quod reliquum est in carne uiuat temporis 4:3 sufficit enim praeteri- 4:3 For the time past of our life may suffice us to have

wrought the will of the Gentiles, when we walked in lasciviousness, lusts, excess of wine, revellings, banquetings,

tum tempus ad uoluntatem gentium consummandam qui am- and abominable idolatries: bulauerunt in luxuriis desideriis uinolentiis comesationibus potationibus et inlicitis idolorum cultibus 4:4 in quo pereg- 4:4 Wherein they think it strange that ye run not with them

to the same excess of riot, speaking evil of you:

rinantur non concurrentibus uobis in eandem luxuriae confusionem blasphemantes 4:5 qui reddent rationem ei qui pa- 4:5 Who shall give account to him that is ready to judge

the quick and the dead.

ratus est iudicare uiuos et mortuos 4:6 propter hoc enim et 4:6 For for this cause was the gospel preached also to them

that are dead, that they might be judged according to men in the flesh, but live according to God in the spirit.

mortuis euangelizatum est ut iudicentur quidem secundum homines in carne uiuant autem secundum Deum spiritu 4:7 4:7 But the end of all things is at hand: be ye therefore

sober, and watch unto prayer.

omnium autem finis adpropinquauit estote itaque prudentes et uigilate in orationibus 4:8 ante omnia mutuam in uosmet 4:8 And above all things have fervent charity among your-

selves: for charity shall cover the multitude of sins.

ipsos caritatem continuam habentes quia caritas operit multitudinem peccatorum 4:9 hospitales inuicem sine murmu- 4:9 Use hospitality one to another without grudging. ratione 4:10 unusquisque sicut accepit gratiam in alterut- 4:10 As every man hath received the gift, even so minister

the same one to another, as good stewards of the manifold grace of God.

rum illam administrantes sicut boni dispensatores multiformis gratiae Dei 4:11 si quis loquitur quasi sermones Dei si 4:11 If any man speak, let him speak as the oracles of God;

if any man minister, let him do it as of the ability which God giveth: that God in all things may be glorified through Jesus Christ, to whom be praise and dominion for ever and ever. Amen.

quis ministrat tamquam ex uirtute quam administrat Deus ut in omnibus honorificetur Deus per Iesum Christum cui est

gloria et imperium in saecula saeculorum amen 4:12 caris- 4:12 Beloved, think it not strange concerning the fiery trial which is to try you, as though some strange thing happened unto you:

simi nolite peregrinari in feruore qui ad temptationem uobis

fit quasi noui aliquid uobis contingat 4:13 sed communi- 4:13 But rejoice, inasmuch as ye are partakers of Christ’s sufferings; that, when his glory shall be revealed, ye may be glad also with exceeding joy.

Biblia Sacra Vulgata

cantes Christi passionibus gaudete ut et in reuelatione glo-

4:14 If ye be reproached for the name of Christ, happy are riae eius gaudeatis exultantes 4:14 si exprobramini in no-

ye; for the spirit of glory and of God resteth upon you: on their part he is evil spoken of, but on your part he is

glorified. mine Christi beati quoniam gloriae Dei Spiritus in uobis re-4:15 But let none of you suffer as a murderer, or as a thief, quiescit 4:15 nemo enim uestrum patiatur quasi homicida

or as an evildoer, or as a busybody in other men’s matters.

4:16 Yet if any man suffer as a Christian, let him not be aut fur aut maledicus aut alienorum appetitor 4:16 si autem

ashamed; but let him glorify God on this behalf.

ut Christianus non erubescat glorificet autem Deum in isto

4:17 For the time is come that judgment must begin at the nomine 4:17 quoniam tempus ut incipiat iudicium de domo

house of God: and if it first begin at us, what shall the end be of them that obey not the gospel of God?

Dei si autem primum a nobis qui finis eorum qui non credunt

4:18 And if the righteous scarcely be saved, where shall Dei euangelio 4:18 et si iustus uix saluatur impius et pec-

the ungodly and the sinner appear?

4:19 Wherefore let them that suffer according to the will cator ubi parebit 4:19 itaque et hii qui patiuntur secundum

of God commit the keeping of their souls to him in well doing, as unto a faithful Creator.

uoluntatem Dei fideli creatori commendant animas suas in benefactis

5:1 The elders which are among you I exhort, who am also 5:1 seniores ergo qui in uobis sunt obsecro consenior et

an elder, and a witness of the sufferings of Christ, and also a partaker of the glory that shall be revealed:

testis Christi passionum qui et eius quae in futuro reue-

5:2 Feed the flock of God which is among you, taking the landa est gloriae communicator 5:2 pascite qui est in uo-

oversight thereof, not by constraint, but willingly; not for filthy lucre, but of a ready mind;

bis gregem Dei prouidentes non coacto sed spontanee se-

5:3 Neither as being lords over God’s heritage, but being cundum Deum neque turpis lucri gratia sed uoluntarie 5:3 ensamples to the flock.

neque ut dominantes in cleris sed formae facti gregi et ex

5:4 And when the chief Shepherd shall appear, ye shall animo 5:4 et cum apparuerit princeps pastorum percipietis

receive a crown of glory that fadeth not away.

5:5 Likewise, ye younger, submit yourselves unto the el- inmarcescibilem gloriae coronam 5:5 similiter adulescen-

der. Yea, all of you be subject one to another, and be clothed with humility: for God resisteth the proud, and

giveth grace to the humble. tes subditi estote senioribus omnes autem inuicem humili-

tatem insinuate quia Deus superbis resistit humilibus autem

5:6 Humble yourselves therefore under the mighty hand dat gratiam 5:6 humiliamini igitur sub potenti manu Dei ut

of God, that he may exalt you in due time:

5:7 Casting all your care upon him; for he careth for you. uos exaltet in tempore uisitationis 5:7 omnem sollicitudi-

nem uestram proicientes in eum quoniam ipsi cura est de

5:8 Be sober, be vigilant; because your adversary the devil, uobis 5:8 sobrii estote uigilate quia aduersarius uester dia-

as a roaring lion, walketh about, seeking whom he may devour:

bolus tamquam leo rugiens circuit quaerens quem deuoret

5:9 Whom resist stedfast in the faith, knowing that the 5:9 cui resistite fortes fide scientes eadem passionum ei quae

same afflictions are accomplished in your brethren that are in the world.

in mundo est uestrae fraternitati fieri 5:10 Deus autem om-

5:10 But the God of all grace, who hath called us unto his eternal glory by Christ Jesus, after that ye have suffered a while, make you perfect, stablish, strengthen, settle you.

nis gratiae qui uocauit nos in aeternam suam gloriam in Christo Iesu modicum passos ipse perficiet confirmabit solida-

5:11 To him be glory and dominion for ever and ever. bit 5:11 ipsi imperium in saecula saeculorum amen 5:12

Amen. 5:12 By Silvanus, a faithful brother unto you, as I suppose, I have written briefly, exhorting, and testifying that this is the true grace of God wherein ye stand.

per Siluanum uobis fidelem fratrem ut arbitror breuiter sc-

Epistola Petri I

ripsi obsecrans et contestans hanc esse ueram gratiam Dei in qua state 5:13 salutat uos quae est in Babylone cumelecta et 5:13 The church that is at Babylon, elected together with

you, saluteth you; and so doth Marcus my son.

Marcus filius meus 5:14 salutate inuicem in osculo sancto 5:14 Greet ye one another with a kiss of charity. Peace be

with you all that are in Christ Jesus. Amen.

gratia uobis omnibus qui estis in Christo

Biblia Sacra Vulgata

MDCCLVIII

Epistola Petri II

1:1 Simon Petrus seruus et apostolus Iesu Christi his qui 1:1 Simon Peter, a servant and an apostle of Jesus Christ,

to them that have obtained like precious faith with us through the righteousness of God and our Saviour Jesus

coaequalem nobis sortiti sunt fidem in iustitia Dei nostri et Christ: saluatoris Iesu Christi 1:2 gratia uobis et pax adimpleatur 1:2 Grace and peace be multiplied unto you through the

knowledge of God, and of Jesus our Lord,

in cognitione Domini nostri 1:3 quomodo omnia nobis di- 1:3 According as his divine power hath given unto us all

things that pertain unto life and godliness, through the knowledge of him that hath called us to glory and virtue:

uinae uirtutis suae quae ad uitam et pietatem donata est per cognitionem eius qui uocauit nos propria gloria et uirtute 1:4 per quae maxima et pretiosa nobis promissa donauit ut 1:4 Whereby are given unto us exceeding great and pre-

cious promises: that by these ye might be partakers of the divine nature, having escaped the corruption that is in the

per haec efficiamini diuinae consortes naturae fugientes eius world through lust.

quae in mundo est concupiscentiae corruptionem 1:5 uos 1:5 And beside this, giving all diligence, add to your faith

virtue; and to virtue knowledge;

autem curam omnem subinferentes ministrate in fide uestra uirtutem in uirtute autem scientiam 1:6 in scientia au- 1:6 And to knowledge temperance; and to temperance pa-

tience; and to patience godliness;

tem abstinentiam in abstinentia autem patientiam in patien-

tia autem pietatem 1:7 in pietate autem amorem fraterni- 1:7 And to godliness brotherly kindness; and to brotherly

kindness charity.

tatis in amore autem fraternitatis caritatem 1:8 haec enim 1:8 For if these things be in you, and abound, they make

you that ye shall neither be barren nor unfruitful in the knowledge of our Lord Jesus Christ.

uobis cum adsint et superent non uacuos nec sine fructu uos constituent in Domini nostri Iesu Christi cognitione 1:9 cui 1:9 But he that lacketh these things is blind, and cannot

see afar off, and hath forgotten that he was purged from his old sins.

enim non praesto sunt haec caecus est et manu temptans obliuionem accipiens purgationis ueterum suorum delictorum 1:10 quapropter fratres magis satagite ut per bona opera cer- 1:10 Wherefore the rather, brethren, give diligence to

make your calling and election sure: for if ye do these things, ye shall never fall:

tam uestram uocationem et electionem faciatis haec enim facientes non peccabitis aliquando 1:11 sic enim abundan- 1:11 For so an entrance shall be ministered unto you

abundantly into the everlasting kingdom of our Lord and Saviour Jesus Christ.

ter ministrabitur uobis introitus in aeternum regnum Domini nostri et saluatoris Iesu Christi 1:12 propter quod incipiam 1:12 Wherefore I will not be negligent to put you always in

remembrance of these things, though ye know them, and be established in the present truth.

uos semper commonere de his et quidem scientes et confirmatos in praesenti ueritate 1:13 iustum autem arbitror 1:13 Yea, I think it meet, as long as I am in this tabernacle,

to stir you up by putting you in remembrance;

quamdiu sum in hoc tabernaculo suscitare uos in commonitione 1:14 certus quod uelox est depositio tabernaculi mei 1:14 Knowing that shortly I must put off this my taberna-

cle, even as our Lord Jesus Christ hath shewed me.

secundum quod et Dominus noster Iesus Christus significauit mihi 1:15 dabo autem operam et frequenter habere uos 1:15 Moreover I will endeavour that ye may be able after

my decease to have these things always in remembrance.

post obitum meum ut horum memoriam faciatis 1:16 non 1:16 For we have not followed cunningly devised fables, when we made known unto you the power and coming of our Lord Jesus Christ, but were eyewitnesses of his

enim doctas fabulas secuti notam fecimus uobis Domini nos- majesty. tri Iesu Christi uirtutem et praesentiam sed speculatores facti illius magnitudinis 1:17 accipiens enim a Deo Patre hono- 1:17 For he received from God the Father honour and

glory, when there came such a voice to him from the excellent glory, This is my beloved Son, in whom I am well pleased.

Biblia Sacra Vulgata

rem et gloriam uoce delapsa ad eum huiuscemodi a magnifica gloria hic est Filius meus dilectus in quo mihi con-

1:18 And this voice which came from heaven we heard, placui 1:18 et hanc uocem nos audiuimus de caelo adlatam

when we were with him in the holy mount.

1:19 We have also a more sure word of prophecy; where- cum essemus cum ipso in monte sancto 1:19 et habemus fir-

unto ye do well that ye take heed, as unto a light that shineth in a dark place, until the day dawn, and the day

star arise in your hearts: miorem propheticum sermonem cui bene facitis adtendentes

quasi lucernae lucenti in caliginoso loco donec dies inluce-

1:20 Knowing this first, that no prophecy of the scripture scat et lucifer oriatur in cordibus uestris 1:20 hoc primum

is of any private interpretation.

intellegentes quod omnis prophetia scripturae propria interp-

1:21 For the prophecy came not in old time by the will of retatione non fit 1:21 non enim uoluntate humana adlata est

man: but holy men of God spake as they were moved by the Holy Ghost.

aliquando prophetia sed Spiritu Sancto inspirati locuti sunt sancti Dei homines

2:1 But there were false prophets also among the peo- 2:1 fuerunt uero et pseudoprophetae in populo sicut et in

ple, even as there shall be false teachers among you, who privily shall bring in damnable heresies, even denying the Lord that bought them, and bring upon themselves swift destruction.

uobis erunt magistri mendaces qui introducent sectas perditionis et eum qui emit eos Dominum negant superducen-

2:2 And many shall follow their pernicious ways; by rea- tes sibi celerem perditionem 2:2 et multi sequentur eorum

son of whom the way of truth shall be evil spoken of.

2:3 And through covetousness shall they with feigned luxurias per quos uia ueritatis blasphemabitur 2:3 et in aua-

words make merchandise of you: whose judgment now of a long time lingereth not, and their damnation slumbereth

not. ritia fictis uerbis de uobis negotiabuntur quibus iudicium iam 2:4 For if God spared not the angels that sinned, but cast olim non cessat et perditio eorum non dormitat 2:4 si enim

them down to hell, and delivered them into chains of darkness, to be reserved unto judgment;

Deus angelis peccantibus non pepercit sed rudentibus inferni detractos in tartarum tradidit in iudicium cruciatos reseruari

2:5 And spared not the old world, but saved Noah the 2:5 et originali mundo non pepercit sed octauum Noe iusti-

eighth person, a preacher of righteousness, bringing in the flood upon the world of the ungodly;

tiae praeconem custodiuit diluuium mundo impiorum indu-

2:6 And turning the cities of Sodom and Gomorrha into cens 2:6 et ciuitates Sodomorum et Gomorraeorum in cine-

ashes condemned them with an overthrow, making them an ensample unto those that after should live ungodly;

rem rediiens euersione damnauit exemplum eorum qui impie

2:7 And delivered just Lot, vexed with the filthy conversa- acturi sunt ponens 2:7 et iustum Loth oppressum a nefantion of the wicked:

2:8 (For that righteous man dwelling among them, in see- dorum iniuria conuersatione eruit 2:8 aspectu enim et au-

ing and hearing, vexed his righteous soul from day to day with their unlawful deeds;)

ditu iustus erat habitans apud eos qui diem de die animam iu-

2:9 The Lord knoweth how to deliver the godly out of stam iniquis operibus cruciabant 2:9 nouit Dominus pios de

temptations, and to reserve the unjust unto the day of judgment to be punished:

temptatione eripere iniquos uero in diem iudicii cruciandos

2:10 But chiefly them that walk after the flesh in the lust reseruare 2:10 magis autem eos qui post carnem in concu-

of uncleanness, and despise government. Presumptuous are they, selfwilled, they are not afraid to speak evil of

dignities. piscentia inmunditiae ambulant dominationemque contem-

nunt audaces sibi placentes sectas non metuunt blaspheman-

2:11 Whereas angels, which are greater in power and tes 2:11 ubi angeli fortitudine et uirtute cum sint maiores

might, bring not railing accusation against them before the Lord.

Epistola Petri II

non portant aduersum se execrabile iudicium 2:12 hii uero 2:12 But these, as natural brute beasts, made to be taken

and destroyed, speak evil of the things that they understand not; and shall utterly perish in their own corruption;

uelut inrationabilia pecora naturaliter in captionem et in perniciem in his quae ignorant blasphemantes in corruptione sua et peribunt 2:13 percipientes mercedem iniustitiae uo- 2:13 And shall receive the reward of unrighteousness, as

they that count it pleasure to riot in the day time. Spots they are and blemishes, sporting themselves with their

luptatem existimantes diei delicias coinquinationes et ma- own deceivings while they feast with you; culae deliciis affluentes in conuiuiis suis luxuriantes uobiscum 2:14 oculos habentes plenos adulterio et incessabiles 2:14 Having eyes full of adultery, and that cannot cease

from sin; beguiling unstable souls: an heart they have exercised with covetous practices; cursed children:

delicti pellicentes animas instabiles cor exercitatum auaritiae habentes maledictionis filii 2:15 derelinquentes rectam 2:15 Which have forsaken the right way, and are gone

astray, following the way of Balaam the son of Bosor, who loved the wages of unrighteousness;

uiam errauerunt secuti uiam Balaam ex Bosor qui mercedem iniquitatis amauit 2:16 correptionem uero habuit suae 2:16 But was rebuked for his iniquity: the dumb ass speak-

ing with man’s voice forbad the madness of the prophet.

uesaniae subiugale mutum in hominis uoce loquens prohibuit prophetae insipientiam 2:17 hii sunt fontes sine aqua et 2:17 These are wells without water, clouds that are carried

with a tempest; to whom the mist of darkness is reserved for ever.

nebulae turbinibus exagitatae quibus caligo tenebrarum reseruatur 2:18 superba enim uanitatis loquentes pellicent in 2:18 For when they speak great swelling words of vanity,

they allure through the lusts of the flesh, through much wantonness, those that were clean escaped from them who

desideriis carnis luxuriae eos qui paululum effugiunt qui in live in error. errore conuersantur 2:19 libertatem illis promittentes cum 2:19 While they promise them liberty, they themselves are

the servants of corruption: for of whom a man is overcome, of the same is he brought in bondage.

ipsi serui sint corruptionis a quo enim quis superatus est huius et seruus est 2:20 si enim refugientes coinquinatio- 2:20 For if after they have escaped the pollutions of the

world through the knowledge of the Lord and Saviour Jesus Christ, they are again entangled therein, and overcome, the latter end is worse with them than the beginning.

nes mundi in cognitione Domini nostri et saluatoris Iesu Christi his rursus inpliciti superantur facta sunt eis posteriora deteriora prioribus 2:21 melius enim erat illis non cognos- 2:21 For it had been better for them not to have known the

way of righteousness, than, after they have known it, to turn from the holy commandment delivered unto them.

cere uiam iustitiae quam post agnitionem retrorsum conuerti ab eo quod illis traditum est sancto mandato 2:22 contigit 2:22 But it is happened unto them according to the true

proverb, The dog is turned to his own vomit again; and the sow that was washed to her wallowing in the mire.

enim eis illud ueri prouerbii canis reuersus ad suum uomitum et sus lota in uolutabro luti

3:1 hanc ecce uobis carissimi secundam scribo epistulam 3:1 This second epistle, beloved, I now write unto you; in

both which I stir up your pure minds by way of remembrance:

in quibus excito uestram in commonitione sinceram mentem 3:2 ut memores sitis eorum quae praedixi uerborum a sanctis 3:2 That ye may be mindful of the words which were spo-

ken before by the holy prophets, and of the commandment of us the apostles of the Lord and Saviour:

prophetis et apostolorum uestrorum praeceptorum Domini et saluatoris 3:3 hoc primum scientes quod uenient in no- 3:3 Knowing this first, that there shall come in the last

days scoffers, walking after their own lusts,

uissimis diebus in deceptione inlusores iuxta proprias concupiscentias ambulantes 3:4 dicentes ubi est promissio aut 3:4 And saying, Where is the promise of his coming? for

since the fathers fell asleep, all things continue as they were from the beginning of the creation.

Biblia Sacra Vulgata

aduentus eius ex quo enim patres dormierunt omnia sic per-

3:5 For this they willingly are ignorant of, that by the word seuerant ab initio creaturae 3:5 latet enim eos hoc uolentes

of God the heavens were of old, and the earth standing out of the water and in the water:

quod caeli erant prius et terra de aqua et per aquam consis-

3:6 Whereby the world that then was, being overflowed tens Dei uerbo 3:6 per quae ille tunc mundus aqua inunda-

with water, perished:

3:7 But the heavens and the earth, which are now, by the tus periit 3:7 caeli autem qui nunc sunt et terra eodem uerbo

same word are kept in store, reserved unto fire against the day of judgment and perdition of ungodly men.

repositi sunt igni seruati in diem iudicii et perditionis im-

3:8 But, beloved, be not ignorant of this one thing, that one piorum hominum 3:8 unum uero hoc non lateat uos carisday is with the Lord as a thousand years, and a thousand years as one day.

simi quia unus dies apud Dominum sicut mille anni et mille

3:9 The Lord is not slack concerning his promise, as some anni sicut dies unus 3:9 non tardat Dominus promissi sed men count slackness; but is longsuffering to us-ward, not willing that any should perish, but that all should come to

repentance. patienter agit propter uos nolens aliquos perire sed omnes 3:10 But the day of the Lord will come as a thief in the ad paenitentiam reuerti 3:10 adueniet autem dies Domini

night; in the which the heavens shall pass away with a great noise, and the elements shall melt with fervent heat, the earth also and the works that are therein shall be burned up.

ut fur in qua caeli magno impetu transient elementa uero ca-

3:11 Seeing then that all these things shall be dissolved, lore soluentur 3:11 cum haec igitur omnia dissoluenda sint

what manner of persons ought ye to be in all holy conversation and godliness,

quales oportet esse uos in sanctis conuersationibus et pieta-

3:12 Looking for and hasting unto the coming of the day of tibus 3:12 expectantes et properantes in aduentum Dei diei

God, wherein the heavens being on fire shall be dissolved, and the elements shall melt with fervent heat?

per quam caeli ardentes soluentur et elementa ignis ardore

3:13 Nevertheless we, according to his promise, look for tabescent 3:13 nouos uero caelos et nouam terram et pronew heavens and a new earth, wherein dwelleth righteousness.

missa ipsius expectamus in quibus iustitia habitat 3:14 pro-

3:14 Wherefore, beloved, seeing that ye look for such things, be diligent that ye may be found of him in peace, without spot, and blameless.

pter quod carissimi haec expectantes satis agite inmaculati et

3:15 And account that the longsuffering of our Lord is sal- inuiolati ei inueniri in pace 3:15 et Domini nostri longani-

vation; even as our beloved brother Paul also according to the wisdom given unto him hath written unto you;

mitatem salutem arbitramini sicut et carissimus frater noster

3:16 As also in all his epistles, speaking in them of these Paulus secundum datam sibi sapientiam scripsit uobis 3:16

things; in which are some things hard to be understood, which they that are unlearned and unstable wrest, as they

do also the other scriptures, unto their own destruction. sicut et in omnibus epistulis loquens in eis de his in quibus sunt quaedam difficilia intellectu quae indocti et instabiles deprauant sicut et ceteras scripturas ad suam ipsorum perdi-

3:17 Ye therefore, beloved, seeing ye know these things tionem 3:17 uos igitur fratres praescientes custodite ne insi-

before, beware lest ye also, being led away with the error of the wicked, fall from your own stedfastness.

pientium errore transducti excidatis a propria firmitate 3:18

3:18 But grow in grace, and in the knowledge of our Lord and Saviour Jesus Christ. To him be glory both now and for ever. Amen.

crescite uero in gratia et in cognitione Domini nostri et saluatoris Iesu Christi ipsi gloria et nunc et in die aeternitatis amen

Epistola Ioannis I

1:1 quod fuit ab initio quod audiuimus quod uidimus ocu- 1:1 That which was from the beginning, which we have

heard, which we have seen with our eyes, which we have looked upon, and our hands have handled, of the Word of

lis nostris quod perspeximus et manus nostrae temptauerunt life; de uerbo uitae 1:2 et uita manifestata est et uidimus et tes- 1:2 (For the life was manifested, and we have seen it, and

bear witness, and shew unto you that eternal life, which was with the Father, and was manifested unto us;)

tamur et adnuntiamus uobis uitam aeternam quae erat apud Patrem et apparuit nobis 1:3 quod uidimus et audiuimus ad- 1:3 That which we have seen and heard declare we unto

you, that ye also may have fellowship with us: and truly our fellowship is with the Father, and with his Son Jesus

nuntiamus et uobis ut et uos societatem habeatis nobiscum Christ. et societas nostra sit cum Patre et cum Filio eius Iesu Christo 1:4 et haec scribimus uobis ut gaudium nostrum sit 1:4 And these things write we unto you, that your joy may

be full.

plenum 1:5 et haec est adnuntiatio quam audiuimus ab eo 1:5 This then is the message which we have heard of him,

and declare unto you, that God is light, and in him is no darkness at all.

et adnuntiamus uobis quoniam Deus lux est et tenebrae in eo non sunt ullae 1:6 si dixerimus quoniam societatem ha- 1:6 If we say that we have fellowship with him, and walk

in darkness, we lie, and do not the truth:

bemus cum eo et in tenebris ambulamus mentimur et non facimus ueritatem 1:7 si autem in luce ambulemus sicut et 1:7 But if we walk in the light, as he is in the light, we

have fellowship one with another, and the blood of Jesus Christ his Son cleanseth us from all sin.

ipse est in luce societatem habemus ad inuicem et sanguis Iesu Filii eius mundat nos ab omni peccato 1:8 si dixeri- 1:8 If we say that we have no sin, we deceive ourselves,

and the truth is not in us.

mus quoniam peccatum non habemus ipsi nos seducimus et ueritas in nobis non est 1:9 si confiteamur peccata nostra fi- 1:9 If we confess our sins, he is faithful and just to forgive

us our sins, and to cleanse us from all unrighteousness.

delis est et iustus ut remittat nobis peccata et emundet nos ab omni iniquitate 1:10 si dixerimus quoniam non peccauimus 1:10 If we say that we have not sinned, we make him a

liar, and his word is not in us.

mendacem facimus eum et uerbum eius non est in nobis

2:1 filioli mei haec scribo uobis ut non peccetis sed et si 2:1 My little children, these things write I unto you, that ye sin not. And if any man sin, we have an advocate with the Father, Jesus Christ the righteous:

quis peccauerit aduocatum habemus apud Patrem Iesum Christum iustum 2:2 et ipse est propitiatio pro peccatis nostris 2:2 And he is the propitiation for our sins: and not for

our’s only, but also for the sins of the whole world.

non pro nostris autem tantum sed etiam pro totius mundi 2:3 et in hoc scimus quoniam cognouimus eum si mandata 2:3 And hereby we do know that we know him, if we keep

his commandments.

eius obseruemus 2:4 qui dicit se nosse eum et mandata eius 2:4 He that saith, I know him, and keepeth not his com-

mandments, is a liar, and the truth is not in him.

non custodit mendax est in hoc ueritas non est 2:5 qui au- 2:5 But whoso keepeth his word, in him verily is the love

of God perfected: hereby know we that we are in him.

tem seruat uerbum eius uere in hoc caritas Dei perfecta est

in hoc scimus quoniam in ipso sumus 2:6 qui dicit se in 2:6 He that saith he abideth in him ought himself also so

to walk, even as he walked.

ipso manere debet sicut ille ambulauit et ipse ambulare 2:7 2:7 Brethren, I write no new commandment unto you, but

an old commandment which ye had from the beginning. The old commandment is the word which ye have heard

carissimi non mandatum nouum scribo uobis sed mandatum from the beginning. uetus quod habuistis ab initio mandatum uetus est uerbum quod audistis 2:8 iterum mandatum nouum scribo uobis 2:8 Again, a new commandment I write unto you, which

thing is true in him and in you: because the darkness is past, and the true light now shineth.

Biblia Sacra Vulgata

quod est uerum et in ipso et in uobis quoniam tenebrae trans-

2:9 He that saith he is in the light, and hateth his brother, eunt et lumen uerum iam lucet 2:9 qui dicit se in luce esse is in darkness even until now. 2:10 He that loveth his brother abideth in the light, and et fratrem suum odit in tenebris est usque adhuc 2:10 qui there is none occasion of stumbling in him.

diligit fratrem suum in lumine manet et scandalum in eo non

2:11 But he that hateth his brother is in darkness, and est 2:11 qui autem odit fratrem suum in tenebris est et in

walketh in darkness, and knoweth not whither he goeth, because that darkness hath blinded his eyes.

tenebris ambulat et nescit quo eat quoniam tenebrae obcae-

2:12 I write unto you, little children, because your sins are cauerunt oculos eius 2:12 scribo uobis filioli quoniam reforgiven you for his name’s sake.

2:13 I write unto you, fathers, because ye have known him mittuntur uobis peccata propter nomen eius 2:13 scribo uo-

that is from the beginning. I write unto you, young men, because ye have overcome the wicked one. I write unto

you, little children, because ye have known the Father. bis patres quoniam cognouistis eum qui ab initio est scribo 2:14 I have written unto you, fathers, because ye have uobis adulescentes quoniam uicistis malignum 2:14 scripsi

known him that is from the beginning. I have written unto you, young men, because ye are strong, and the word of

God abideth in you, and ye have overcome the wicked one. uobis infantes quoniam cognouistis Patrem scripsi uobis patres quia cognouistis eum qui ab initio scripsi uobis adulescentes quia fortes estis et uerbum Dei in uobis manet et uici-

2:15 Love not the world, neither the things that are in the stis malignum 2:15 nolite diligere mundum neque ea quae world. If any man love the world, the love of the Father is not in him.

in mundo sunt si quis diligit mundum non est caritas Patris in

2:16 For all that is in the world, the lust of the flesh, and the eo 2:16 quoniam omne quod est in mundo concupiscentia lust of the eyes, and the pride of life, is not of the Father, but is of the world.

carnis et concupiscentia oculorum est et superbia uitae quae

2:17 And the world passeth away, and the lust thereof: but non est ex Patre sed ex mundo est 2:17 et mundus transit he that doeth the will of God abideth for ever.

et concupiscentia eius qui autem facit uoluntatem Dei manet

2:18 Little children, it is the last time: and as ye have heard in aeternum 2:18 filioli nouissima hora est et sicut audistis that antichrist shall come, even now are there many antichrists; whereby we know that it is the last time.

quia antichristus uenit nunc antichristi multi facti sunt unde

2:19 They went out from us, but they were not of us; for scimus quoniam nouissima hora est 2:19 ex nobis prodii-

if they had been of us, they would no doubt have continued with us: but they went out, that they might be made

manifest that they were not all of us. runt sed non erant ex nobis nam si fuissent ex nobis per-

mansissent utique nobiscum sed ut manifesti sint quoniam

2:20 But ye have an unction from the Holy One, and ye non sunt omnes ex nobis 2:20 sed uos unctionem habetis

know all things.

2:21 I have not written unto you because ye know not the a Sancto et nostis omnia 2:21 non scripsi uobis quasi ig-

truth, but because ye know it, and that no lie is of the truth.

norantibus ueritatem sed quasi scientibus eam et quoniam

2:22 Who is a liar but he that denieth that Jesus is the omne mendacium ex ueritate non est 2:22 quis est mendax

Christ? He is antichrist, that denieth the Father and the Son.

nisi is qui negat quoniam Iesus non est Christus hic est an-

2:23 Whosoever denieth the Son, the same hath not the Fa- tichristus qui negat Patrem et Filium 2:23 omnis qui negat

ther: he that acknowledgeth the Son hath the Father also.

Filium nec Patrem habet qui confitetur Filium et Patrem ha-

2:24 Let that therefore abide in you, which ye have heard bet 2:24 uos quod audistis ab initio in uobis permaneat si

from the beginning. If that which ye have heard from the beginning shall remain in you, ye also shall continue in

the Son, and in the Father. in uobis permanserit quod ab initio audistis et uos in Filio et

Epistola Ioannis I

Patre manebitis 2:25 et haec est repromissio quam ipse pol- 2:25 And this is the promise that he hath promised us, even

eternal life.

licitus est nobis uitam aeternam 2:26 haec scripsi uobis de 2:26 These things have I written unto you concerning them

that seduce you.

eis qui seducunt uos 2:27 et uos unctionem quam accepistis 2:27 But the anointing which ye have received of him

abideth in you, and ye need not that any man teach you: but as the same anointing teacheth you of all things, and is truth, and is no lie, and even as it hath taught you, ye shall abide in him.

ab eo manet in uobis et non necesse habetis ut aliquis doceat uos sed sicut unctio eius docet uos de omnibus et uerum est

et non est mendacium et sicut docuit uos manete in eo 2:28 2:28 And now, little children, abide in him; that, when he shall appear, we may have confidence, and not be ashamed before him at his coming.

et nunc filioli manete in eo ut cum apparuerit habeamus fiduciam et non confundamur ab eo in aduentu eius 2:29 si 2:29 If ye know that he is righteous, ye know that every

one that doeth righteousness is born of him.

scitis quoniam iustus est scitote quoniam et omnis qui facit iustitiam ex ipso natus est

3:1 uidete qualem caritatem dedit nobis Pater ut filii Dei 3:1 Behold, what manner of love the Father hath bestowed

upon us, that we should be called the sons of God: therefore the world knoweth us not, because it knew him not.

nominemur et sumus propter hoc mundus non nouit nos quia non nouit eum 3:2 carissimi nunc filii Dei sumus et non- 3:2 Beloved, now are we the sons of God, and it doth not

yet appear what we shall be: but we know that, when he shall appear, we shall be like him; for we shall see him as

dum apparuit quid erimus scimus quoniam cum apparuerit he is. similes ei erimus quoniam uidebimus eum sicuti est 3:3 et 3:3 And every man that hath this hope in him purifieth

himself, even as he is pure.

omnis qui habet spem hanc in eo sanctificat se sicut et ille sanctus est 3:4 omnis qui facit peccatum et iniquitatem fa- 3:4 Whosoever committeth sin transgresseth also the law:

for sin is the transgression of the law.

cit et peccatum est iniquitas 3:5 et scitis quoniam ille appa- 3:5 And ye know that he was manifested to take away our

sins; and in him is no sin.

ruit ut peccata tolleret et peccatum in eo non est 3:6 omnis 3:6 Whosoever abideth in him sinneth not: whosoever sin-

neth hath not seen him, neither known him.

qui in eo manet non peccat omnis qui peccat non uidit eum nec cognouit eum 3:7 filioli nemo uos seducat qui facit iu- 3:7 Little children, let no man deceive you: he that doeth

righteousness is righteous, even as he is righteous.

stitiam iustus est sicut et ille iustus est 3:8 qui facit pecca- 3:8 He that committeth sin is of the devil; for the devil

sinneth from the beginning. For this purpose the Son of God was manifested, that he might destroy the works of

tum ex diabolo est quoniam ab initio diabolus peccat in hoc the devil. apparuit Filius Dei ut dissoluat opera diaboli 3:9 omnis qui 3:9 Whosoever is born of God doth not commit sin; for

his seed remaineth in him: and he cannot sin, because he is born of God.

natus est ex Deo peccatum non facit quoniam semen ipsius in eo manet et non potest peccare quoniam ex Deo natus est 3:10 in hoc manifesti sunt filii Dei et filii diaboli omnis qui 3:10 In this the children of God are manifest, and the chil-

dren of the devil: whosoever doeth not righteousness is not of God, neither he that loveth not his brother.

non est iustus non est de Deo et qui non diligit fratrem suum 3:11 quoniam haec est adnuntiatio quam audistis ab initio ut 3:11 For this is the message that ye heard from the begin-

ning, that we should love one another.

diligamus alterutrum 3:12 non sicut Cain ex maligno erat 3:12 Not as Cain, who was of that wicked one, and slew

his brother. And wherefore slew he him? Because his own works were evil, and his brother’s righteous.

et occidit fratrem suum et propter quid occidit eum quoniam opera eius maligna erant fratris autem eius iusta 3:13 nolite 3:13 Marvel not, my brethren, if the world hate you. mirari fratres si odit uos mundus 3:14 nos scimus quoniam 3:14 We know that we have passed from death unto life,

because we love the brethren. He that loveth not his brother abideth in death.

Biblia Sacra Vulgata

translati sumus de morte in uitam quoniam diligimus fratres

3:15 Whosoever hateth his brother is a murderer: and ye qui non diligit manet in morte 3:15 omnis qui odit fratrem

know that no murderer hath eternal life abiding in him.

suum homicida est et scitis quoniam omnis homicida non

3:16 Hereby perceive we the love of God, because he laid habet uitam aeternam in se manentem 3:16 in hoc cognoui-

down his life for us: and we ought to lay down our lives for the brethren.

mus caritatem quoniam ille pro nobis animam suam posuit et

3:17 But whoso hath this world’s good, and seeth his nos debemus pro fratribus animas ponere 3:17 qui habuerit

brother have need, and shutteth up his bowels of compassion from him, how dwelleth the love of God in him?

substantiam mundi et uiderit fratrem suum necesse habere et clauserit uiscera sua ab eo quomodo caritas Dei manet in eo

3:18 My little children, let us not love in word, neither in 3:18 filioli non diligamus uerbo nec lingua sed opere et uetongue; but in deed and in truth.

3:19 And hereby we know that we are of the truth, and ritate 3:19 in hoc cognoscimus quoniam ex ueritate sumus

shall assure our hearts before him.

3:20 For if our heart condemn us, God is greater than our et in conspectu eius suadeamus corda nostra 3:20 quoniam

heart, and knoweth all things.

si reprehenderit nos cor maior est Deus corde nostro et nouit

3:21 Beloved, if our heart condemn us not, then have we omnia 3:21 carissimi si cor non reprehenderit nos fiduciam

confidence toward God.

3:22 And whatsoever we ask, we receive of him, because habemus ad Deum 3:22 et quodcumque petierimus accipie-

we keep his commandments, and do those things that are pleasing in his sight.

mus ab eo quoniam mandata eius custodimus et ea quae sunt

3:23 And this is his commandment, That we should be- placita coram eo facimus 3:23 et hoc est mandatum eius ut

lieve on the name of his Son Jesus Christ, and love one another, as he gave us commandment.

credamus in nomine Filii eius Iesu Christi et diligamus alte-

3:24 And he that keepeth his commandments dwelleth in rutrum sicut dedit mandatum nobis 3:24 et qui seruat man-

him, and he in him. And hereby we know that he abideth in us, by the Spirit which he hath given us.

data eius in illo manet et ipse in eo et in hoc scimus quoniam manet in nobis de Spiritu quem nobis dedit

4:1 Beloved, believe not every spirit, but try the spirits 4:1 carissimi nolite omni spiritui credere sed probate spi-

whether they are of God: because many false prophets are gone out into the world.

ritus si ex Deo sint quoniam multi pseudoprophetae exierunt

4:2 Hereby know ye the Spirit of God: Every spirit that in mundum 4:2 in hoc cognoscitur Spiritus Dei omnis spi-

confesseth that Jesus Christ is come in the flesh is of God:

ritus qui confitetur Iesum Christum in carne uenisse ex Deo

4:3 And every spirit that confesseth not that Jesus Christ est 4:3 et omnis spiritus qui soluit Iesum ex Deo non est is come in the flesh is not of God: and this is that spirit of antichrist, whereof ye have heard that it should come; and

even now already is it in the world. et hoc est antichristi quod audistis quoniam uenit et nunc 4:4 Ye are of God, little children, and have overcome iam in mundo est 4:4 uos ex Deo estis filioli et uicistis eos

them: because greater is he that is in you, than he that is in the world.

quoniam maior est qui in uobis est quam qui in mundo 4:5

4:5 They are of the world: therefore speak they of the world, and the world heareth them.

ipsi de mundo sunt ideo de mundo loquuntur et mundus eos

4:6 We are of God: he that knoweth God heareth us; he audit 4:6 nos ex Deo sumus qui nouit Deum audit nos qui

that is not of God heareth not us. Hereby know we the spirit of truth, and the spirit of error.

non est ex Deo non audit nos in hoc cognoscimus Spiritum

4:7 Beloved, let us love one another: for love is of God; ueritatis et spiritum erroris 4:7 carissimi diligamus inuicem

and every one that loveth is born of God, and knoweth God.

quoniam caritas ex Deo est et omnis qui diligit ex Deo natus

Epistola Ioannis I

est et cognoscit Deum 4:8 qui non diligit non nouit Deum 4:8 He that loveth not knoweth not God; for God is love. quoniam Deus caritas est 4:9 in hoc apparuit caritas Dei in 4:9 In this was manifested the love of God toward us, be-

cause that God sent his only begotten Son into the world, that we might live through him.

nobis quoniam Filium suum unigenitum misit Deus in mundum ut uiuamus per eum 4:10 in hoc est caritas non quasi 4:10 Herein is love, not that we loved God, but that he

loved us, and sent his Son to be the propitiation for our sins.

nos dilexerimus Deum sed quoniam ipse dilexit nos et misit Filium suum propitiationem pro peccatis nostris 4:11 caris- 4:11 Beloved, if God so loved us, we ought also to love

one another.

simi si sic Deus dilexit nos et nos debemus alterutrum diligere 4:12 Deum nemo uidit umquam si diligamus inui- 4:12 No man hath seen God at any time. If we love one

another, God dwelleth in us, and his love is perfected in us.

cem Deus in nobis manet et caritas eius in nobis perfecta est 4:13 in hoc intellegimus quoniam in eo manemus et ipse in 4:13 Hereby know we that we dwell in him, and he in us,

because he hath given us of his Spirit.

nobis quoniam de Spiritu suo dedit nobis 4:14 et nos uidi- 4:14 And we have seen and do testify that the Father sent

the Son to be the Saviour of the world.

mus et testificamur quoniam Pater misit Filium saluatorem mundi 4:15 quisque confessus fuerit quoniam Iesus est Fi- 4:15 Whosoever shall confess that Jesus is the Son of God,

God dwelleth in him, and he in God.

lius Dei Deus in eo manet et ipse in Deo 4:16 et nos cogno- 4:16 And we have known and believed the love that God

hath to us. God is love; and he that dwelleth in love dwelleth in God, and God in him.

uimus et credidimus caritati quam habet Deus in nobis Deus caritas est et qui manet in caritate in Deo manet et Deus in eo 4:17 in hoc perfecta est caritas nobiscum ut fiduciam ha- 4:17 Herein is our love made perfect, that we may have

boldness in the day of judgment: because as he is, so are we in this world.

beamus in die iudicii quia sicut ille est et nos sumus in hoc mundo 4:18 timor non est in caritate sed perfecta caritas fo- 4:18 There is no fear in love; but perfect love casteth out

fear: because fear hath torment. He that feareth is not made perfect in love.

ras mittit timorem quoniam timor poenam habet qui autem timet non est perfectus in caritate 4:19 nos ergo diligamus 4:19 We love him, because he first loved us. quoniam Deus prior dilexit nos 4:20 si quis dixerit quoniam 4:20 If a man say, I love God, and hateth his brother, he

is a liar: for he that loveth not his brother whom he hath seen, how can he love God whom he hath not seen?

diligo Deum et fratrem suum oderit mendax est qui enim non diligit fratrem suum quem uidit Deum quem non uidit quomodo potest diligere 4:21 et hoc mandatum habemus ab eo 4:21 And this commandment have we from him, That he

who loveth God love his brother also.

ut qui diligit Deum diligat et fratrem suum

5:1 omnis qui credit quoniam Iesus est Christus ex Deo 5:1 Whosoever believeth that Jesus is the Christ is born of God: and every one that loveth him that begat loveth him also that is begotten of him.

natus est et omnis qui diligit eum qui genuit diligit eum qui natus est ex eo 5:2 in hoc cognoscimus quoniam diligimus 5:2 By this we know that we love the children of God,

when we love God, and keep his commandments.

natos Dei cum Deum diligamus et mandata eius faciamus 5:3 haec est enim caritas Dei ut mandata eius custodiamus 5:3 For this is the love of God, that we keep his command-

ments: and his commandments are not grievous.

et mandata eius grauia non sunt 5:4 quoniam omne quod 5:4 For whatsoever is born of God overcometh the world: and this is the victory that overcometh the world, even our faith.

natum est ex Deo uincit mundum et haec est uictoria quae

Biblia Sacra Vulgata

5:5 Who is he that overcometh the world, but he that be- uincit mundum fides nostra 5:5 quis est qui uincit mundum

lieveth that Jesus is the Son of God?

5:6 This is he that came by water and blood, even Jesus nisi qui credit quoniam Iesus est Filius Dei 5:6 hic est qui

Christ; not by water only, but by water and blood. And it is the Spirit that beareth witness, because the Spirit is

truth. uenit per aquam et sanguinem Iesus Christus non in aqua

solum sed in aqua et sanguine et Spiritus est qui testificatur

5:7 For there are three that bear record in heaven, the Fa- quoniam Christus est ueritas 5:7 quia tres sunt qui testimo-

ther, the Word, and the Holy Ghost: and these three are one.

nium dant 5:8 Spiritus et aqua et sanguis et tres unum sunt

5:8 And there are three that bear witness in earth, the Spirit, and the water, and the blood: and these three agree in one.

5:9 si testimonium hominum accipimus testimonium Dei ma-

5:9 If we receive the witness of men, the witness of God is greater: for this is the witness of God which he hath

testified of his Son. ius est quoniam hoc est testimonium Dei quod maius est quia 5:10 He that believeth on the Son of God hath the witness testificatus est de Filio suo 5:10 qui credit in Filio Dei habet

in himself: he that believeth not God hath made him a liar; because he believeth not the record that God gave of his

Son. testimonium Dei in se qui non credit Filio mendacem facit

eum quoniam non credidit in testimonio quod testificatus est

5:11 And this is the record, that God hath given to us eter- Deus de Filio suo 5:11 et hoc est testimonium quoniam uinal life, and this life is in his Son.

tam aeternam dedit nobis Deus et haec uita in Filio eius est

5:12 He that hath the Son hath life; and he that hath not 5:12 qui habet Filium habet uitam qui non habet Filium Dei

the Son of God hath not life.

5:13 These things have I written unto you that believe on uitam non habet 5:13 haec scripsi uobis ut sciatis quoniam

the name of the Son of God; that ye may know that ye have eternal life, and that ye may believe on the name of

the Son of God. uitam habetis aeternam qui creditis in nomine Filii Dei 5:14

5:14 And this is the confidence that we have in him, that, if we ask any thing according to his will, he heareth us:

et haec est fiducia quam habemus ad eum quia quodcum-

5:15 And if we know that he hear us, whatsoever we ask, que petierimus secundum uoluntatem eius audit nos 5:15 et

we know that we have the petitions that we desired of him.

scimus quoniam audit nos quicquid petierimus scimus quo-

5:16 If any man see his brother sin a sin which is not unto niam habemus petitiones quas postulauimus ab eo 5:16 qui

death, he shall ask, and he shall give him life for them that sin not unto death. There is a sin unto death: I do not say

that he shall pray for it. scit fratrem suum peccare peccatum non ad mortem petet

et dabit ei uitam peccantibus non ad mortem est peccatum

5:17 All unrighteousness is sin: and there is a sin not unto ad mortem non pro illo dico ut roget 5:17 omnis iniquitas death. 5:18 We know that whosoever is born of God sinneth not; peccatum est et est peccatum non ad mortem 5:18 scimus but he that is begotten of God keepeth himself, and that wicked one toucheth him not.

quoniam omnis qui natus est ex Deo non peccat sed gene-

5:19 And we know that we are of God, and the whole ratio Dei conseruat eum et malignus non tangit eum 5:19

world lieth in wickedness.

scimus quoniam ex Deo sumus et mundus totus in maligno

5:20 And we know that the Son of God is come, and hath positus est 5:20 et scimus quoniam Filius Dei uenit et de-

given us an understanding, that we may know him that is true, and we are in him that is true, even in his Son Jesus

Christ. This is the true God, and eternal life. dit nobis sensum ut cognoscamus uerum Deum et simus in 5:21 Little children, keep yourselves from idols. Amen. uero Filio eius hic est uerus Deus et uita aeterna 5:21 filioli

custodite uos a simulacris

Epistola Ioannis II

1:1 senior electae dominae et natis eius quos ego diligo 1:1 The elder unto the elect lady and her children, whom I

love in the truth; and not I only, but also all they that have known the truth;

in ueritate et non ego solus sed et omnes qui cognouerunt ueritatem 1:2 propter ueritatem quae permanet in nobis et 1:2 For the truth’s sake, which dwelleth in us, and shall be

with us for ever.

nobiscum erit in aeternum 1:3 sit nobiscum gratia miseri- 1:3 Grace be with you, mercy, and peace, from God the

Father, and from the Lord Jesus Christ, the Son of the Father, in truth and love.

cordia pax a Deo Patre et a Christo Iesu Filio Patris in ueritate et caritate 1:4 gauisus sum ualde quoniam inueni de 1:4 I rejoiced greatly that I found of thy children walking

in truth, as we have received a commandment from the Father.

filiis tuis ambulantes in ueritate sicut mandatum accepimus a

Patre 1:5 et nunc rogo te domina non tamquam mandatum 1:5 And now I beseech thee, lady, not as though I wrote a new commandment unto thee, but that which we had from the beginning, that we love one another.

nouum scribens tibi sed quod habuimus ab initio ut diligamus alterutrum 1:6 et haec est caritas ut ambulemus secun- 1:6 And this is love, that we walk after his command-

ments. This is the commandment, That, as ye have heard from the beginning, ye should walk in it.

dum mandata eius hoc mandatum est ut quemadmodum audistis ab initio in eo ambuletis 1:7 quoniam multi seducto- 1:7 For many deceivers are entered into the world, who

confess not that Jesus Christ is come in the flesh. This is a deceiver and an antichrist.

res exierunt in mundum qui non confitentur Iesum Christum uenientem in carne hic est seductor et antichristus 1:8 ui- 1:8 Look to yourselves, that we lose not those things

which we have wrought, but that we receive a full reward.

dete uosmet ipsos ne perdatis quae operati estis sed ut merce-

dem plenam accipiatis 1:9 omnis qui praecedit et non ma- 1:9 Whosoever transgresseth, and abideth not in the doctrine of Christ, hath not God. He that abideth in the doctrine of Christ, he hath both the Father and the Son.

net in doctrina Christi Deum non habet qui permanet in doctrina hic et Filium et Patrem habet 1:10 si quis uenit ad uos 1:10 If there come any unto you, and bring not this doc-

trine, receive him not into your house, neither bid him God speed:

et hanc doctrinam non adfert nolite recipere eum in domum nec haue ei dixeritis 1:11 qui enim dicit illi haue commu- 1:11 For he that biddeth him God speed is partaker of his

evil deeds.

nicat operibus illius malignis 1:12 plura habens uobis scri- 1:12 Having many things to write unto you, I would not

write with paper and ink: but I trust to come unto you, and speak face to face, that our joy may be full.

bere nolui per cartam et atramentum spero enim me futurum apud uos et os ad os loqui ut gaudium uestrum plenum sit

1:13 salutant te filii sororis tuae electae 1:13 The children of thy elect sister greet thee. Amen.

Biblia Sacra Vulgata

MDCCLXX

Epistola Ioannis III

1:1 senior Gaio carissimo quem ego diligo in ueritate 1:2 1:1 The elder unto the wellbeloved Gaius, whom I love in

the truth.

1:2 Beloved, I wish above all things that thou mayest prosper and be in health, even as thy soul prospereth.

carissime de omnibus orationem facio prospere te ingredi et ualere sicut prospere agit anima tua 1:3 gauisus sum ualde 1:3 For I rejoiced greatly, when the brethren came and tes-

tified of the truth that is in thee, even as thou walkest in the truth.

uenientibus fratribus et testimonium perhibentibus ueritati

tuae sicut tu in ueritate ambulas 1:4 maiorem horum non 1:4 I have no greater joy than to hear that my children walk

in truth.

habeo gratiam quam ut audiam filios meos in ueritate ambulantes 1:5 carissime fideliter facis quicquid operaris in fra- 1:5 Beloved, thou doest faithfully whatsoever thou doest

to the brethren, and to strangers;

tres et hoc in peregrinos 1:6 qui testimonium reddiderunt 1:6 Which have borne witness of thy charity before the

church: whom if thou bring forward on their journey after a godly sort, thou shalt do well:

caritati tuae in conspectu ecclesiae quos bene facies deducens digne Deo 1:7 pro nomine enim profecti sunt nihil ac- 1:7 Because that for his name’s sake they went forth, tak-

ing nothing of the Gentiles.

cipientes a gentibus 1:8 nos ergo debemus suscipere huius- 1:8 We therefore ought to receive such, that we might be

fellowhelpers to the truth.

modi ut cooperatores simus ueritatis 1:9 scripsissem forsi- 1:9 I wrote unto the church: but Diotrephes, who loveth to

have the preeminence among them, receiveth us not.

tan ecclesiae sed is qui amat primatum gerere in eis Diotrepes non recipit nos 1:10 propter hoc si uenero commo- 1:10 Wherefore, if I come, I will remember his deeds

which he doeth, prating against us with malicious words: and not content therewith, neither doth he himself receive the brethren, and forbiddeth them that would, and casteth them out of the church.

neam eius opera quae facit uerbis malignis garriens in nos et quasi non ei ista sufficiant nec ipse suscipit fratres et eos qui cupiunt prohibet et de ecclesia eicit 1:11 carissime noli 1:11 Beloved, follow not that which is evil, but that which

is good. He that doeth good is of God: but he that doeth evil hath not seen God.

imitari malum sed quod bonum est qui benefacit ex Deo est qui malefacit non uidit Deum 1:12 Demetrio testimonium 1:12 Demetrius hath good report of all men, and of the

truth itself: yea, and we also bear record; and ye know that our record is true.

redditur ab omnibus et ab ipsa ueritate et nos autem testimonium perhibemus et nosti quoniam testimonium nostrum uerum est 1:13 multa habui scribere tibi sed nolui per atra- 1:13 I had many things to write, but I will not with ink and

pen write unto thee:

mentum et calamum scribere tibi 1:14 spero autem protinus 1:14 But I trust I shall shortly see thee, and we shall speak

face to face. Peace be to thee. Our friends salute thee. Greet the friends by name.

te uidere et os ad os loquemur 1:15 pax tibi salutant te amici

1:15 Peace be to thee. Our friends salute thee. Greet the friends by name.

saluta amicos per nomen

Biblia Sacra Vulgata

MDCCLXXII

Epistola Catholica Iudae

1:1 Iudas Iesu Christi seruus frater autem Iacobi his qui 1:1 Jude, the servant of Jesus Christ, and brother of James, to them that are sanctified by God the Father, and preserved in Jesus Christ, and called:

in Deo Patre dilectis et Iesu Christo conseruatis uocatis 1:2

1:2 Mercy unto you, and peace, and love, be multiplied.

misericordia uobis et pax et caritas adimpleatur 1:3 caris- 1:3 Beloved, when I gave all diligence to write unto you

of the common salvation, it was needful for me to write unto you, and exhort you that ye should earnestly contend

simi omnem sollicitudinem faciens scribendi uobis de com- for the faith which was once delivered unto the saints. muni uestra salute necesse habui scribere uobis deprecans supercertari semel traditae sanctis fidei 1:4 subintroierunt 1:4 For there are certain men crept in unawares, who were

before of old ordained to this condemnation, ungodly men, turning the grace of our God into lasciviousness, and deny-

enim quidam homines qui olim praescripti sunt in hoc iu- ing the only Lord God, and our Lord Jesus Christ. dicium impii Dei nostri gratiam transferentes in luxuriam et solum Dominatorem et Dominum nostrum Iesum Christum negantes 1:5 commonere autem uos uolo scientes se- 1:5 I will therefore put you in remembrance, though ye

once knew this, how that the Lord, having saved the people out of the land of Egypt, afterward destroyed them that

mel omnia quoniam Iesus populum de terra Aegypti sal- believed not.

uans secundo eos qui non crediderunt perdidit 1:6 angelos 1:6 And the angels which kept not their first estate, but left their own habitation, he hath reserved in everlasting chains under darkness unto the judgment of the great day.

uero qui non seruauerunt suum principatum sed dereliquerunt suum domicilium in iudicium magni diei uinculis aeternis sub caligine reseruauit 1:7 sicut Sodoma et Gomorra 1:7 Even as Sodom and Gomorrha, and the cities about

them in like manner, giving themselves over to fornication, and going after strange flesh, are set forth for an ex-

et finitimae ciuitates simili modo exfornicatae et abeuntes ample, suffering the vengeance of eternal fire. post carnem alteram factae sunt exemplum ignis aeterni poenam sustinentes 1:8 similiter et hii carnem quidem macu- 1:8 Likewise also these filthy dreamers defile the flesh,

despise dominion, and speak evil of dignities.

lant dominationem autem spernunt maiestates autem blas-

phemant 1:9 cum Michahel archangelus cum diabolo dis- 1:9 Yet Michael the archangel, when contending with the devil he disputed about the body of Moses, durst not bring against him a railing accusation, but said, The Lord rebuke

putans altercaretur de Mosi corpore non est ausus iudicium thee. inferre blasphemiae sed dixit imperet tibi Dominus 1:10 hii 1:10 But these speak evil of those things which they know

not: but what they know naturally, as brute beasts, in those things they corrupt themselves.

autem quaecumque quidem ignorant blasphemant quaecumque autem naturaliter tamquam muta animalia norunt in his corrumpuntur 1:11 uae illis quia uia Cain abierunt et er- 1:11 Woe unto them! for they have gone in the way of

Cain, and ran greedily after the error of Balaam for reward, and perished in the gainsaying of Core.

rore Balaam mercede effusi sunt et contradictione Core perierunt 1:12 hii sunt in epulis suis maculae conuiuantes sine 1:12 These are spots in your feasts of charity, when they

feast with you, feeding themselves without fear: clouds they are without water, carried about of winds; trees whose fruit withereth, without fruit, twice dead, plucked up by the roots;

timore semet ipsos pascentes nubes sine aqua quae a uentis circumferuntur arbores autumnales infructuosae bis mortuae eradicatae 1:13 fluctus feri maris despumantes suas confu- 1:13 Raging waves of the sea, foaming out their own

shame; wandering stars, to whom is reserved the blackness of darkness for ever.

siones sidera errantia quibus procella tenebrarum in aeternum seruata est 1:14 prophetauit autem et his septimus ab 1:14 And Enoch also, the seventh from Adam, prophesied

of these, saying, Behold, the Lord cometh with ten thousands of his saints,

Adam Enoc dicens ecce uenit Dominus in sanctis milibus

Biblia Sacra Vulgata

1:15 To execute judgment upon all, and to convince all suis 1:15 facere iudicium contra omnes et arguere omnes

that are ungodly among them of all their ungodly deeds which they have ungodly committed, and of all their hard

speeches which ungodly sinners have spoken against him. impios de omnibus operibus impietatis eorum quibus impie

egerunt et de omnibus duris quae locuti sunt contra eum pec-

1:16 These are murmurers, complainers, walking after catores impii 1:16 hii sunt murmuratores querellosi secun-

their own lusts; and their mouth speaketh great swelling words, having men’s persons in admiration because of ad-

vantage. dum desideria sua ambulantes et os illorum loquitur superba 1:17 But, beloved, remember ye the words which were mirantes personas quaestus causa 1:17 uos autem carissimi

spoken before of the apostles of our Lord Jesus Christ;

memores estote uerborum quae praedicta sunt ab apostolis

1:18 How that they told you there should be mockers in the Domini nostri Iesu Christi 1:18 quia dicebant uobis quo-

last time, who should walk after their own ungodly lusts.

niam in nouissimo tempore uenient inlusores secundum sua

1:19 These be they who separate themselves, sensual, hav- desideria ambulantes impietatum 1:19 hii sunt qui segreing not the Spirit. 1:20 But ye, beloved, building up yourselves on your most gant animales Spiritum non habentes 1:20 uos autem caholy faith, praying in the Holy Ghost,

rissimi superaedificantes uosmet ipsos sanctissimae uestrae

1:21 Keep yourselves in the love of God, looking for the fidei in Spiritu Sancto orantes 1:21 ipsos uos in dilectione

mercy of our Lord Jesus Christ unto eternal life.

1:22 And of some have compassion, making a difference: Dei seruate 1:22 et hos quidem arguite iudicatos 1:23 il-

1:23 And others save with fear, pulling them out of the fire; hating even the garment spotted by the flesh.

los uero saluate de igne rapientes aliis autem miseremini in timore odientes et eam quae carnalis est maculatam tunicam

1:24 Now unto him that is able to keep you from falling, 1:24 ei autem qui potest uos conseruare sine peccato et con-

and to present you faultless before the presence of his glory with exceeding joy,

stituere ante conspectum gloriae suae inmaculatos in exul-

1:25 To the only wise God our Saviour, be glory and tatione 1:25 soli Deo saluatori nostro per Iesum Christum

majesty, dominion and power, both now and ever. Amen.

Dominum nostrum gloria magnificentia imperium et potestas ante omne saeculum et nunc et in omnia saecula amen

Apocalypsis Ionannis

1:1 apocalypsis Iesu Christi quam dedit illi Deus palam fa- 1:1 The Revelation of Jesus Christ, which God gave unto

him, to shew unto his servants things which must shortly come to pass; and he sent and signified it by his angel unto

cere seruis suis quae oportet fieri cito et significauit mittens his servant John: per angelum suum seruo suo Iohanni 1:2 qui testimonium 1:2 Who bare record of the word of God, and of the testi-

mony of Jesus Christ, and of all things that he saw.

perhibuit uerbo Dei et testimonium Iesu Christi quaecumque uidit 1:3 beatus qui legit et qui audiunt uerba prophe- 1:3 Blessed is he that readeth, and they that hear the words

of this prophecy, and keep those things which are written therein: for the time is at hand.

tiae et seruant ea quae in ea scripta sunt tempus enim prope est 1:4 Iohannes septem ecclesiis quae sunt in Asia gratia 1:4 John to the seven churches which are in Asia: Grace be

unto you, and peace, from him which is, and which was, and which is to come; and from the seven Spirits which

uobis et pax ab eo qui est et qui erat et qui uenturus est et are before his throne; a septem spiritibus qui in conspectu throni eius sunt 1:5 et 1:5 And from Jesus Christ, who is the faithful witness, and

the first begotten of the dead, and the prince of the kings of the earth. Unto him that loved us, and washed us from

ab Iesu Christo qui est testis fidelis primogenitus mortuo- our sins in his own blood, rum et princeps regum terrae qui dilexit nos et lauit nos a peccatis nostris in sanguine suo 1:6 et fecit nostrum reg- 1:6 And hath made us kings and priests unto God and his

Father; to him be glory and dominion for ever and ever. Amen.

num sacerdotes Deo et Patri suo ipsi gloria et imperium in saecula saeculorum amen 1:7 ecce uenit cum nubibus et ui- 1:7 Behold, he cometh with clouds; and every eye shall see

him, and they also which pierced him: and all kindreds of the earth shall wail because of him. Even so, Amen.

debit eum omnis oculus et qui eum pupugerunt et plangent se super eum omnes tribus terrae etiam amen 1:8 ego sum 1:8 I am Alpha and Omega, the beginning and the ending,

saith the Lord, which is, and which was, and which is to come, the Almighty.

Alpha et Omega principium et finis dicit Dominus Deus qui est et qui erat et qui uenturus est Omnipotens 1:9 ego Io- 1:9 I John, who also am your brother, and companion

in tribulation, and in the kingdom and patience of Jesus Christ, was in the isle that is called Patmos, for the word

hannes frater uester et particeps in tribulatione et regno et of God, and for the testimony of Jesus Christ. patientia in Iesu fui in insula quae appellatur Patmos propter uerbum Dei et testimonium Iesu 1:10 fui in spiritu in 1:10 I was in the Spirit on the Lord’s day, and heard behind

me a great voice, as of a trumpet,

dominica die et audiui post me uocem magnam tamquam tubae 1:11 dicentis quod uides scribe in libro et mitte septem 1:11 Saying, I am Alpha and Omega, the first and the last:

and, What thou seest, write in a book, and send it unto the seven churches which are in Asia; unto Ephesus, and unto Smyrna, and unto Pergamos, and unto Thyatira, and unto Sardis, and unto Philadelphia, and unto Laodicea.

ecclesiis Ephesum et Zmyrnam et Pergamum et Thyatiram et Sardis et Philadelphiam et Laodiciam 1:12 et conuersus 1:12 And I turned to see the voice that spake with me. And

being turned, I saw seven golden candlesticks;

sum ut uiderem uocem quae loquebatur mecum et conuersus uidi septem candelabra aurea 1:13 et in medio septem can- 1:13 And in the midst of the seven candlesticks one like

unto the Son of man, clothed with a garment down to the foot, and girt about the paps with a golden girdle.

delabrorum similem Filio hominis uestitum podere et prae-

cinctum ad mamillas zonam auream 1:14 caput autem eius 1:14 His head and his hairs were white like wool, as white

as snow; and his eyes were as a flame of fire;

et capilli erant candidi tamquam lana alba tamquam nix et oculi eius uelut flamma ignis 1:15 et pedes eius similes ori- 1:15 And his feet like unto fine brass, as if they burned in

a furnace; and his voice as the sound of many waters.

chalco sicut in camino ardenti et uox illius tamquam uox aquarum multarum 1:16 et habebat in dextera sua stellas 1:16 And he had in his right hand seven stars: and out of

his mouth went a sharp twoedged sword: and his countenance was as the sun shineth in his strength.

Biblia Sacra Vulgata

septem et de ore eius gladius utraque parte acutus exiebat et

1:17 And when I saw him, I fell at his feet as dead. And facies eius sicut sol lucet in uirtute sua 1:17 et cum uidis-

he laid his right hand upon me, saying unto me, Fear not; I am the first and the last:

sem eum cecidi ad pedes eius tamquam mortuus et posuit dexteram suam super me dicens noli timere ego sum primus

1:18 I am he that liveth, and was dead; and, behold, I am et nouissimus 1:18 et uiuus et fui mortuus et ecce sum ui-

alive for evermore, Amen; and have the keys of hell and of death.

uens in saecula saeculorum et habeo claues mortis et inferni

1:19 Write the things which thou hast seen, and the things 1:19 scribe ergo quae uidisti et quae sunt et quae oportet fieri

which are, and the things which shall be hereafter;

1:20 The mystery of the seven stars which thou sawest in post haec 1:20 sacramentum septem stellarum quas uidisti

my right hand, and the seven golden candlesticks. The seven stars are the angels of the seven churches: and the seven candlesticks which thou sawest are the seven churches.

in dextera mea et septem candelabra aurea septem stellae angeli sunt septem ecclesiarum et candelabra septem septem ecclesiae sunt

2:1 Unto the angel of the church of Ephesus write; These 2:1 angelo Ephesi ecclesiae scribe haec dicit qui tenet sep-

things saith he that holdeth the seven stars in his right hand, who walketh in the midst of the seven golden can-

dlesticks; tem stellas in dextera sua qui ambulat in medio septem can-2:2 I know thy works, and thy labour, and thy patience, delabrorum aureorum 2:2 scio opera tua et laborem et pa-

and how thou canst not bear them which are evil: and thou hast tried them which say they are apostles, and are not,

and hast found them liars: tientiam tuam et quia non potes sustinere malos et temptasti

eos qui se dicunt apostolos et non sunt et inuenisti eos men-

2:3 And hast borne, and hast patience, and for my name’s daces 2:3 et patientiam habes et sustinuisti propter nomen

sake hast laboured, and hast not fainted. 2:4 Nevertheless I have somewhat against thee, because meum et non defecisti 2:4 sed habeo aduersus te quod cathou hast left thy first love.

2:5 Remember therefore from whence thou art fallen, and ritatem tuam primam reliquisti 2:5 memor esto itaque unde

repent, and do the first works; or else I will come unto thee quickly, and will remove thy candlestick out of his place,

except thou repent. excideris et age paenitentiam et prima opera fac sin autem

uenio tibi et mouebo candelabrum tuum de loco suo nisi pae-

2:6 But this thou hast, that thou hatest the deeds of the nitentiam egeris 2:6 sed hoc habes quia odisti facta Nico-

Nicolaitanes, which I also hate. 2:7 He that hath an ear, let him hear what the Spirit saith laitarum quae et ego odi 2:7 qui habet aurem audiat quid unto the churches; To him that overcometh will I give to eat of the tree of life, which is in the midst of the paradise

of God. Spiritus dicat ecclesiis uincenti dabo ei edere de ligno uitae 2:8 And unto the angel of the church in Smyrna write; quod est in paradiso Dei mei 2:8 et angelo Zmyrnae eccle-

These things saith the first and the last, which was dead, and is alive;

siae scribe haec dicit primus et nouissimus qui fuit mortuus

2:9 I know thy works, and tribulation, and poverty, (but et uiuit 2:9 scio tribulationem tuam et paupertatem tuam

thou art rich) and I know the blasphemy of them which say they are Jews, and are not, but are the synagogue of

Satan. sed diues es et blasphemaris ab his qui se dicunt Iudaeos 2:10 Fear none of those things which thou shalt suffer: esse et non sunt sed sunt synagoga Satanae 2:10 nihil ho-

behold, the devil shall cast some of you into prison, that ye may be tried; and ye shall have tribulation ten days: be thou faithful unto death, and I will give thee a crown of life.

rum timeas quae passurus es ecce missurus est diabolus ex uobis in carcerem ut temptemini et habebitis tribulationem diebus decem esto fidelis usque ad mortem et dabo tibi co-

2:11 He that hath an ear, let him hear what the Spirit saith ronam uitae 2:11 qui habet aurem audiat quid Spiritus dicat

unto the churches; He that overcometh shall not be hurt of the second death.

Apocalypsis Ionannis

ecclesiis qui uicerit non laedetur a morte secunda 2:12 et 2:12 And to the angel of the church in Pergamos write;

These things saith he which hath the sharp sword with two edges;

angelo Pergami ecclesiae scribe haec dicit qui habet rompheam utraque parte acutam 2:13 scio ubi habitas ubi se- 2:13 I know thy works, and where thou dwellest, even

where Satan’s seat is: and thou holdest fast my name, and hast not denied my faith, even in those days wherein Antipas was my faithful martyr, who was slain among you, where Satan dwelleth.

des est Satanae et tenes nomen meum et non negasti fidem meam et in diebus Antipas testis meus fidelis qui occisus est apud uos ubi Satanas habitat 2:14 sed habeo aduersus 2:14 But I have a few things against thee, because thou

hast there them that hold the doctrine of Balaam, who taught Balac to cast a stumblingblock before the children of Israel, to eat things sacrificed unto idols, and to commit fornication.

te pauca quia habes illic tenentes doctrinam Balaam qui docebat Balac mittere scandalum coram filiis Israhel edere et fornicari 2:15 ita habes et tu tenentes doctrinam Nicolaita- 2:15 So hast thou also them that hold the doctrine of the

Nicolaitanes, which thing I hate.

rum 2:16 similiter paenitentiam age si quo minus uenio tibi 2:16 Repent; or else I will come unto thee quickly, and

will fight against them with the sword of my mouth.

cito et pugnabo cum illis in gladio oris mei 2:17 qui habet 2:17 He that hath an ear, let him hear what the Spirit saith

unto the churches; To him that overcometh will I give to eat of the hidden manna, and will give him a white stone, and in the stone a new name written, which no man knoweth saving he that receiveth it.

aurem audiat quid Spiritus dicat ecclesiis uincenti dabo ei manna absconditum et dabo illi calculum candidum et in calculo nomen nouum scriptum quod nemo scit nisi qui accipit 2:18 et angelo Thyatirae ecclesiae scribe haec dicit Filius Dei 2:18 And unto the angel of the church in Thyatira write;

These things saith the Son of God, who hath his eyes like unto a flame of fire, and his feet are like fine brass;

qui habet oculos ut flammam ignis et pedes eius similes orichalco 2:19 noui opera tua et caritatem et fidem et ministe- 2:19 I know thy works, and charity, and service, and faith,

and thy patience, and thy works; and the last to be more than the first.

rium et patientiam tuam et opera tua nouissima plura prio-

ribus 2:20 sed habeo aduersus te quia permittis mulierem 2:20 Notwithstanding I have a few things against thee, because thou sufferest that woman Jezebel, which calleth herself a prophetess, to teach and to seduce my servants to

Hiezabel quae se dicit propheten docere et seducere seruos commit fornication, and to eat things sacrificed unto idols.

meos fornicari et manducare de idolothytis 2:21 et dedi illi 2:21 And I gave her space to repent of her fornication; and

she repented not.

tempus ut paenitentiam ageret et non uult paeniteri a fornicatione sua 2:22 ecce mitto eam in lectum et qui moechantur 2:22 Behold, I will cast her into a bed, and them that com-

mit adultery with her into great tribulation, except they repent of their deeds.

cum ea in tribulationem maximam nisi paenitentiam egerint ab operibus eius 2:23 et filios eius interficiam in morte et 2:23 And I will kill her children with death; and all the

churches shall know that I am he which searcheth the reins and hearts: and I will give unto every one of you according

scient omnes ecclesiae quia ego sum scrutans renes et corda to your works. et dabo unicuique uestrum secundum opera uestra 2:24 uo- 2:24 But unto you I say, and unto the rest in Thyatira, as

many as have not this doctrine, and which have not known the depths of Satan, as they speak; I will put upon you

bis autem dico ceteris qui Thyatirae estis quicumque non none other burden. habent doctrinam hanc qui non cognouerunt altitudines Satanae quemadmodum dicunt non mittam super uos aliud pon-

dus 2:25 tamen id quod habetis tenete donec ueniam 2:26 2:25 But that which ye have already hold fast till I come. 2:26 And he that overcometh, and keepeth my works unto the end, to him will I give power over the nations:

et qui uicerit et qui custodierit usque in finem opera mea dabo illi potestatem super gentes 2:27 et reget illas in uirga 2:27 And he shall rule them with a rod of iron; as the ves-

sels of a potter shall they be broken to shivers: even as I received of my Father.

Biblia Sacra Vulgata

2:28 And I will give him the morning star. ferrea tamquam uas figuli confringentur 2:28 sicut et ego

2:29 He that hath an ear, let him hear what the Spirit saith accepi a Patre meo et dabo illi stellam matutinam 2:29 qui unto the churches.

habet aurem audiat quid Spiritus dicat ecclesiis

3:1 And unto the angel of the church in Sardis write; These 3:1 et angelo ecclesiae Sardis scribe haec dicit qui habet

things saith he that hath the seven Spirits of God, and the seven stars; I know thy works, that thou hast a name that

thou livest, and art dead. septem spiritus Dei et septem stellas scio opera tua quia no-

3:2 Be watchful, and strengthen the things which remain, men habes quod uiuas et mortuus es 3:2 esto uigilans et that are ready to die: for I have not found thy works perfect before God.

confirma cetera quae moritura erant non enim inuenio opera

3:3 Remember therefore how thou hast received and tua plena coram Deo meo 3:3 in mente ergo habe qualiter

heard, and hold fast, and repent. If therefore thou shalt not watch, I will come on thee as a thief, and thou shalt

not know what hour I will come upon thee. acceperis et audieris et serua et paenitentiam age si ergo non

uigilaueris ueniam tamquam fur et nescies qua hora ueniam

3:4 Thou hast a few names even in Sardis which have ad te 3:4 sed habes pauca nomina in Sardis qui non inqui-

not defiled their garments; and they shall walk with me in white: for they are worthy.

nauerunt uestimenta sua et ambulabunt mecum in albis quia

3:5 He that overcometh, the same shall be clothed in white digni sunt 3:5 qui uicerit sic uestietur uestimentis albis et

raiment; and I will not blot out his name out of the book of life, but I will confess his name before my Father, and

before his angels. non delebo nomen eius de libro uitae et confitebor nomen

3:6 He that hath an ear, let him hear what the Spirit saith eius coram Patre meo et coram angelis eius 3:6 qui habet unto the churches.

3:7 And to the angel of the church in Philadelphia write; aurem audiat quid Spiritus dicat ecclesiis 3:7 et angelo Phi-

These things saith he that is holy, he that is true, he that hath the key of David, he that openeth, and no man shut-

teth; and shutteth, and no man openeth; ladelphiae ecclesiae scribe haec dicit sanctus et uerus qui ha-

bet clauem Dauid qui aperit et nemo cludit et cludit et nemo

3:8 I know thy works: behold, I have set before thee an aperit 3:8 scio opera tua ecce dedi coram te ostium aper-

open door, and no man can shut it: for thou hast a little strength, and hast kept my word, and hast not denied my

name. tum quod nemo potest cludere quia modicam habes uirtutem

3:9 Behold, I will make them of the synagogue of Satan, et seruasti uerbum meum et non negasti nomen meum 3:9 which say they are Jews, and are not, but do lie; behold, I will make them to come and worship before thy feet, and

to know that I have loved thee. ecce dabo de synagoga Satanae qui dicunt se Iudaeos esse et

non sunt sed mentiuntur ecce faciam illos ut ueniant et ad-

3:10 Because thou hast kept the word of my patience, I orent ante pedes tuos et scient quia ego dilexi te 3:10 quo-

also will keep thee from the hour of temptation, which shall come upon all the world, to try them that dwell upon

the earth. niam seruasti uerbum patientiae meae et ego te seruabo ab

hora temptationis quae uentura est in orbem uniuersum tem-

3:11 Behold, I come quickly: hold that fast which thou ptare habitantes in terra 3:11 uenio cito tene quod habes ut

hast, that no man take thy crown.

3:12 Him that overcometh will I make a pillar in the tem- nemo accipiat coronam tuam 3:12 qui uicerit faciam illum

ple of my God, and he shall go no more out: and I will write upon him the name of my God, and the name of the city of my God, which is new Jerusalem, which cometh down out of heaven from my God: and I will write upon him my new name.

columnam in templo Dei mei et foras non egredietur amplius et scribam super eum nomen Dei mei et nomen ciuitatis Dei mei nouae Hierusalem quae descendit de caelo a Deo meo

3:13 He that hath an ear, let him hear what the Spirit saith et nomen meum nouum 3:13 qui habet aurem audiat quid unto the churches.

3:14 And unto the angel of the church of the Laodiceans Spiritus dicat ecclesiis 3:14 et angelo Laodiciae ecclesiae

write; These things saith the Amen, the faithful and true witness, the beginning of the creation of God;

Apocalypsis Ionannis

scribe haec dicit Amen testis fidelis et uerus qui est principium creaturae Dei 3:15 scio opera tua quia neque frigidus 3:15 I know thy works, that thou art neither cold nor hot:

I would thou wert cold or hot.

es neque calidus utinam frigidus esses aut calidus 3:16 sed 3:16 So then because thou art lukewarm, and neither cold

nor hot, I will spue thee out of my mouth.

quia tepidus es et nec frigidus nec calidus incipiam te euomere ex ore meo 3:17 quia dicis quod diues sum et locuple- 3:17 Because thou sayest, I am rich, and increased with

goods, and have need of nothing; and knowest not that thou art wretched, and miserable, and poor, and blind, and

tatus et nullius egeo et nescis quia tu es miser et miserabilis naked:

et pauper et caecus et nudus 3:18 suadeo tibi emere a me 3:18 I counsel thee to buy of me gold tried in the fire, that thou mayest be rich; and white raiment, that thou mayest be clothed, and that the shame of thy nakedness do not appear; and anoint thine eyes with eyesalve, that thou mayest see.

aurum ignitum probatum ut locuples fias et uestimentis albis induaris et non appareat confusio nuditatis tuae et collyrio inungue oculos tuos ut uideas 3:19 ego quos amo arguo et 3:19 As many as I love, I rebuke and chasten: be zealous

therefore, and repent.

castigo aemulare ergo et paenitentiam age 3:20 ecce sto ad 3:20 Behold, I stand at the door, and knock: if any man

hear my voice, and open the door, I will come in to him, and will sup with him, and he with me.

ostium et pulso si quis audierit uocem meam et aperuerit ianuam introibo ad illum et cenabo cum illo et ipse mecum 3:21 qui uicerit dabo ei sedere mecum in throno meo sicut et 3:21 To him that overcometh will I grant to sit with me in

my throne, even as I also overcame, and am set down with my Father in his throne.

ego uici et sedi cum Patre meo in throno eius 3:22 qui habet

3:22 He that hath an ear, let him hear what the Spirit saith unto the churches.

aurem audiat quid Spiritus dicat ecclesiis

4:1 post haec uidi et ecce ostium apertum in caelo et uox 4:1 After this I looked, and, behold, a door was opened in heaven: and the first voice which I heard was as it were of a trumpet talking with me; which said, Come up hither,

prima quam audiui tamquam tubae loquentis mecum dicens and I will shew thee things which must be hereafter. ascende huc et ostendam tibi quae oportet fieri post haec 4:2 statim fui in spiritu et ecce sedis posita erat in caelo et 4:2 And immediately I was in the spirit: and, behold, a

throne was set in heaven, and one sat on the throne.

supra sedem sedens 4:3 et qui sedebat similis erat aspectui 4:3 And he that sat was to look upon like a jasper and a

sardine stone: and there was a rainbow round about the throne, in sight like unto an emerald.

lapidis iaspidis et sardini et iris erat in circuitu sedis similis uisioni zmaragdinae 4:4 et in circuitu sedis sedilia uiginti 4:4 And round about the throne were four and twenty

seats: and upon the seats I saw four and twenty elders sitting, clothed in white raiment; and they had on their heads

quattuor et super thronos uiginti quattuor seniores sedentes crowns of gold. circumamictos uestimentis albis et in capitibus eorum coronas aureas 4:5 et de throno procedunt fulgura et uoces 4:5 And out of the throne proceeded lightnings and thun-

derings and voices: and there were seven lamps of fire burning before the throne, which are the seven Spirits of

et tonitrua et septem lampades ardentes ante thronum quae God. sunt septem spiritus Dei 4:6 et in conspectu sedis tamquam 4:6 And before the throne there was a sea of glass like unto

crystal: and in the midst of the throne, and round about the throne, were four beasts full of eyes before and behind.

mare uitreum simile cristallo et in medio sedis et in circuitu sedis quattuor animalia plena oculis ante et retro 4:7 et ani- 4:7 And the first beast was like a lion, and the second beast

like a calf, and the third beast had a face as a man, and the fourth beast was like a flying eagle.

mal primum simile leoni et secundum animal simile uitulo et tertium animal habens faciem quasi hominis et quartum animal simile aquilae uolanti 4:8 et quattuor animalia singula 4:8 And the four beasts had each of them six wings about

him; and they were full of eyes within: and they rest not day and night, saying, Holy, holy, holy, LORD God Almighty, which was, and is, and is to come.

Biblia Sacra Vulgata

eorum habebant alas senas et in circuitu et intus plena sunt oculis et requiem non habent die et nocte dicentia sanctus sanctus sanctus Dominus Deus omnipotens qui erat et qui

4:9 And when those beasts give glory and honour and est et qui uenturus est 4:9 et cum darent illa animalia glo-

thanks to him that sat on the throne, who liveth for ever and ever,

riam et honorem et benedictionem sedenti super thronum ui-

4:10 The four and twenty elders fall down before him that uenti in saecula saeculorum 4:10 procident uiginti quattuor

sat on the throne, and worship him that liveth for ever and ever, and cast their crowns before the throne, saying,

seniores ante sedentem in throno et adorabunt uiuentem in saecula saeculorum et mittent coronas suas ante thronum di-

4:11 Thou art worthy, O Lord, to receive glory and honour centes 4:11 dignus es Domine et Deus noster accipere glo-

and power: for thou hast created all things, and for thy pleasure they are and were created.

riam et honorem et uirtutem quia tu creasti omnia et propter uoluntatem tuam erant et creata sunt

5:1 And I saw in the right hand of him that sat on the 5:1 et uidi in dextera sedentis super thronum librum sc-

throne a book written within and on the backside, sealed with seven seals.

riptum intus et foris signatum sigillis septem 5:2 et uidi

5:2 And I saw a strong angel proclaiming with a loud voice, Who is worthy to open the book, and to loose the seals thereof?

angelum fortem praedicantem uoce magna quis est dignus

5:3 And no man in heaven, nor in earth, neither under the aperire librum et soluere signacula eius 5:3 et nemo pote-

earth, was able to open the book, neither to look thereon.

rat in caelo neque in terra neque subtus terram aperire lib-

5:4 And I wept much, because no man was found worthy rum neque respicere illum 5:4 et ego flebam multum quo-

to open and to read the book, neither to look thereon.

niam nemo dignus inuentus est aperire librum nec uidere

5:5 And one of the elders saith unto me, Weep not: be- eum 5:5 et unus de senioribus dicit mihi ne fleueris ecce

hold, the Lion of the tribe of Juda, the Root of David, hath prevailed to open the book, and to loose the seven seals

thereof. uicit leo de tribu Iuda radix Dauid aperire librum et septem 5:6 And I beheld, and, lo, in the midst of the throne and signacula eius 5:6 et uidi et ecce in medio throni et quattuor

of the four beasts, and in the midst of the elders, stood a Lamb as it had been slain, having seven horns and seven eyes, which are the seven Spirits of God sent forth into all the earth.

animalium et in medio seniorum agnum stantem tamquam occisum habentem cornua septem et oculos septem qui sunt

5:7 And he came and took the book out of the right hand spiritus Dei missi in omnem terram 5:7 et uenit et accepit

of him that sat upon the throne.

5:8 And when he had taken the book, the four beasts and de dextera sedentis de throno 5:8 et cum aperuisset librum

four and twenty elders fell down before the Lamb, having every one of them harps, and golden vials full of odours,

which are the prayers of saints. quattuor animalia et uiginti quattuor seniores ceciderunt co-

ram agno habentes singuli citharas et fialas aureas plenas

5:9 And they sung a new song, saying, Thou art worthy odoramentorum quae sunt orationes sanctorum 5:9 et can-

to take the book, and to open the seals thereof: for thou wast slain, and hast redeemed us to God by thy blood out

of every kindred, and tongue, and people, and nation; tant nouum canticum dicentes dignus es accipere librum et aperire signacula eius quoniam occisus es et redeeisti nos Deo in sanguine tuo ex omni tribu et lingua et populo et na-

5:10 And hast made us unto our God kings and priests: tione 5:10 et fecisti eos Deo nostro regnum et sacerdotes et

and we shall reign on the earth.

5:11 And I beheld, and I heard the voice of many angels regnabunt super terram 5:11 et uidi et audiui uocem ange-

round about the throne and the beasts and the elders: and the number of them was ten thousand times ten thousand, and thousands of thousands;

Apocalypsis Ionannis

lorum multorum in circuitu throni et animalium et seniorum et erat numerus eorum milia milium 5:12 dicentium uoce 5:12 Saying with a loud voice, Worthy is the Lamb that

was slain to receive power, and riches, and wisdom, and strength, and honour, and glory, and blessing.

magna dignus est agnus qui occisus est accipere uirtutem et diuinitatem et sapientiam et fortitudinem et honorem et gloriam et benedictionem 5:13 et omnem creaturam quae in 5:13 And every creature which is in heaven, and on the

earth, and under the earth, and such as are in the sea, and all that are in them, heard I saying, Blessing, and honour, and glory, and power, be unto him that sitteth upon the throne, and unto the Lamb for ever and ever.

caelo est et super terram et sub terram et quae sunt in mari et quae in ea omnes audiui dicentes sedenti in throno et agno benedictio et honor et gloria et potestas in saecula saeculorum 5:14 et quattuor animalia dicebant amen et seniores 5:14 And the four beasts said, Amen. And the four and

twenty elders fell down and worshipped him that liveth for ever and ever.

ceciderunt et adorauerunt

6:1 et uidi quod aperuisset agnus unum de septem signa- 6:1 And I saw when the Lamb opened one of the seals,

and I heard, as it were the noise of thunder, one of the four beasts saying, Come and see.

culis et audiui unum de quattuor animalibus dicentem tamquam uocem tonitrui ueni 6:2 et uidi et ecce equus albus 6:2 And I saw, and behold a white horse: and he that sat

on him had a bow; and a crown was given unto him: and he went forth conquering, and to conquer.

et qui sedebat super illum habebat arcum et data est ei corona et exiuit uincens ut uinceret 6:3 et cum aperuisset si- 6:3 And when he had opened the second seal, I heard the

second beast say, Come and see.

gillum secundum audiui secundum animal dicens ueni 6:4 6:4 And there went out another horse that was red: and

power was given to him that sat thereon to take peace from the earth, and that they should kill one another: and there

et exiuit alius equus rufus et qui sedebat super illum datum was given unto him a great sword. est ei ut sumeret pacem de terra et ut inuicem se interficiant et datus est illi gladius magnus 6:5 et cum aperuisset si- 6:5 And when he had opened the third seal, I heard the

third beast say, Come and see. And I beheld, and lo a black horse; and he that sat on him had a pair of balances

gillum tertium audiui tertium animal dicens ueni et uidi et in his hand. ecce equus niger et qui sedebat super eum habebat stateram

in manu sua 6:6 et audiui tamquam uocem in medio quat- 6:6 And I heard a voice in the midst of the four beasts say, A measure of wheat for a penny, and three measures of barley for a penny; and see thou hurt not the oil and the

tuor animalium dicentem bilibris tritici denario et tres bilib- wine. res hordei denario et uinum et oleum ne laeseris 6:7 et cum 6:7 And when he had opened the fourth seal, I heard the

voice of the fourth beast say, Come and see.

aperuisset sigillum quartum audiui uocem quarti animalis dicentis ueni et uidi 6:8 et ecce equus pallidus et qui sedebat 6:8 And I looked, and behold a pale horse: and his name

that sat on him was Death, and Hell followed with him. And power was given unto them over the fourth part of the earth, to kill with sword, and with hunger, and with death, and with the beasts of the earth.

desuper nomen illi Mors et inferus sequebatur eum et data est illi potestas super quattuor partes terrae interficere gladio fame et morte et bestiis terrae 6:9 et cum aperuisset quin- 6:9 And when he had opened the fifth seal, I saw under the

altar the souls of them that were slain for the word of God, and for the testimony which they held:

tum sigillum uidi subtus altare animas interfectorum propter uerbum Dei et propter testimonium quod habebant 6:10 et 6:10 And they cried with a loud voice, saying, How long,

O Lord, holy and true, dost thou not judge and avenge our blood on them that dwell on the earth?

clamabant uoce magna dicentes usquequo Domine sanctus et uerus non iudicas et uindicas sanguinem nostrum de his

Biblia Sacra Vulgata

6:11 And white robes were given unto every one of them; qui habitant in terra 6:11 et datae sunt illis singulae stolae

and it was said unto them, that they should rest yet for a little season, until their fellowservants also and their brethren, that should be killed as they were, should be fulfilled.

albae et dictum est illis ut requiescerent tempus adhuc modicum donec impleantur conserui eorum et fratres eorum qui

6:12 And I beheld when he had opened the sixth seal, and, interficiendi sunt sicut et illi 6:12 et uidi cum aperuisset si-

lo, there was a great earthquake; and the sun became black as sackcloth of hair, and the moon became as blood;

gillum sextum et terraemotus factus est magnus et sol factus est niger tamquam saccus cilicinus et luna tota facta est sic-

6:13 And the stars of heaven fell unto the earth, even as a ut sanguis 6:13 et stellae caeli ceciderunt super terram sicut

fig tree casteth her untimely figs, when she is shaken of a mighty wind.

ficus mittit grossos suos cum uento magno mouetur 6:14 et

6:14 And the heaven departed as a scroll when it is rolled together; and every mountain and island were moved out of their places.

caelum recessit sicut liber inuolutus et omnis mons et insu-

6:15 And the kings of the earth, and the great men, and the lae de locis suis motae sunt 6:15 et reges terrae et principes

rich men, and the chief captains, and the mighty men, and every bondman, and every free man, hid themselves in the

dens and in the rocks of the mountains; et tribuni et diuites et fortes et omnis seruus et liber abs-6:16 And said to the mountains and rocks, Fall on us, and conderunt se in speluncis et petris montium 6:16 et dicunt

hide us from the face of him that sitteth on the throne, and from the wrath of the Lamb:

montibus et petris cadite super nos et abscondite nos a facie

6:17 For the great day of his wrath is come; and who shall sedentis super thronum et ab ira agni 6:17 quoniam uenit

be able to stand?

dies magnus irae ipsorum et quis poterit stare

7:1 And after these things I saw four angels standing on 7:1 post haec uidi quattuor angelos stantes super quattuor

the four corners of the earth, holding the four winds of the earth, that the wind should not blow on the earth, nor on

the sea, nor on any tree. angulos terrae tenentes quattuor uentos terrae ne flaret uen-

tus super terram neque super mare neque in ullam arborem

7:2 And I saw another angel ascending from the east, hav- 7:2 et uidi alterum angelum ascendentem ab ortu solis ha-

ing the seal of the living God: and he cried with a loud voice to the four angels, to whom it was given to hurt the

earth and the sea, bentem signum Dei uiui et clamauit uoce magna quattuor 7:3 Saying, Hurt not the earth, neither the sea, nor the angelis quibus datum est nocere terrae et mari 7:3 dicens

trees, till we have sealed the servants of our God in their foreheads.

nolite nocere terrae neque mari neque arboribus quoadusque

7:4 And I heard the number of them which were sealed: signemus seruos Dei nostri in frontibus eorum 7:4 et audiui

and there were sealed an hundred and forty and four thousand of all the tribes of the children of Israel.

numerum signatorum centum quadraginta quattuor milia si-

7:5 Of the tribe of Juda were sealed twelve thousand. Of gnati ex omni tribu filiorum Israhel 7:5 ex tribu Iuda duo-

the tribe of Reuben were sealed twelve thousand. Of the tribe of Gad were sealed twelve thousand.

decim milia signati ex tribu Ruben duodecim milia ex tribu

7:6 Of the tribe of Aser were sealed twelve thousand. Of Gad duodecim milia 7:6 ex tribu Aser duodecim milia ex the tribe of Nephthalim were sealed twelve thousand. Of the tribe of Manasses were sealed twelve thousand.

tribu Nepthalim duodecim milia ex tribu Manasse duodecim

7:7 Of the tribe of Simeon were sealed twelve thousand. milia 7:7 ex tribu Symeon duodecim milia ex tribu Leui Of the tribe of Levi were sealed twelve thousand. Of the tribe of Issachar were sealed twelve thousand.

duodecim milia ex tribu Issachar duodecim milia 7:8 ex

7:8 Of the tribe of Zabulon were sealed twelve thousand. Of the tribe of Joseph were sealed twelve thousand. Of the tribe of Benjamin were sealed twelve thousand.

tribu Zabulon duodecim milia ex tribu Ioseph duodecim mi-

7:9 After this I beheld, and, lo, a great multitude, which lia ex tribu Beniamin duodecim milia signati 7:9 post haec

no man could number, of all nations, and kindreds, and people, and tongues, stood before the throne, and before the Lamb, clothed with white robes, and palms in their hands;

uidi turbam magnam quam dinumerare nemo poterat ex om-

Apocalypsis Ionannis

nibus gentibus et tribubus et populis et linguis stantes ante thronum et in conspectu agni amicti stolas albas et palmae

in manibus eorum 7:10 et clamabant uoce magna dicentes 7:10 And cried with a loud voice, saying, Salvation to our

God which sitteth upon the throne, and unto the Lamb.

salus Deo nostro qui sedet super thronum et agno 7:11 et 7:11 And all the angels stood round about the throne, and about the elders and the four beasts, and fell before the throne on their faces, and worshipped God,

omnes angeli stabant in circuitu throni et seniorum et quattuor animalium et ceciderunt in conspectu throni in facies suas et adorauerunt Deum 7:12 dicentes amen benedictio 7:12 Saying, Amen: Blessing, and glory, and wisdom, and

thanksgiving, and honour, and power, and might, be unto our God for ever and ever. Amen.

et claritas et sapientia et gratiarum actio et honor et uirtus et fortitudo Deo nostro in saecula saeculorum amen 7:13 7:13 And one of the elders answered, saying unto me,

What are these which are arrayed in white robes? and whence came they?

et respondit unus de senioribus dicens mihi hii qui amicti sunt stolis albis qui sunt et unde uenerunt 7:14 et dixi illi 7:14 And I said unto him, Sir, thou knowest. And he said

to me, These are they which came out of great tribulation, and have washed their robes, and made them white in the

domine mi tu scis et dixit mihi hii sunt qui ueniunt de tribu- blood of the Lamb. latione magna et lauerunt stolas suas et dealbauerunt eas in sanguine agni 7:15 ideo sunt ante thronum Dei et seruiunt 7:15 Therefore are they before the throne of God, and

serve him day and night in his temple: and he that sitteth on the throne shall dwell among them.

ei die ac nocte in templo eius et qui sedet in throno habitabit super illos 7:16 non esurient neque sitient amplius neque 7:16 They shall hunger no more, neither thirst any more;

neither shall the sun light on them, nor any heat.

cadet super illos sol neque ullus aestus 7:17 quoniam ag- 7:17 For the Lamb which is in the midst of the throne shall

feed them, and shall lead them unto living fountains of waters: and God shall wipe away all tears from their eyes.

nus qui in medio throni est reget illos et deducet eos ad uitae fontes aquarum et absterget Deus omnem lacrimam ex oculis eorum

8:1 et cum aperuisset sigillum septimum factum est silen- 8:1 And when he had opened the seventh seal, there was

silence in heaven about the space of half an hour.

tium in caelo quasi media hora 8:2 et uidi septem angelos 8:2 And I saw the seven angels which stood before God;

and to them were given seven trumpets.

stantes in conspectu Dei et datae sunt illis septem tubae 8:3 8:3 And another angel came and stood at the altar, having

a golden censer; and there was given unto him much incense, that he should offer it with the prayers of all saints

et alius angelus uenit et stetit ante altare habens turibulum upon the golden altar which was before the throne. aureum et data sunt illi incensa multa ut daret orationibus sanctorum omnium super altare aureum quod est ante thronum 8:4 et ascendit fumus incensorum de orationibus sanc- 8:4 And the smoke of the incense, which came with the

prayers of the saints, ascended up before God out of the angel’s hand.

torum de manu angeli coram Deo 8:5 et accepit angelus tu-

8:5 And the angel took the censer, and filled it with fire of the altar, and cast it into the earth: and there were voices, and thunderings, and lightnings, and an earthquake.

ribulum et impleuit illud de igne altaris et misit in terram et facta sunt tonitrua et uoces et fulgora et terraemotus 8:6 8:6 And the seven angels which had the seven trumpets

prepared themselves to sound.

et septem angeli qui habebant septem tubas parauerunt se ut tuba canerent 8:7 et primus tuba cecinit et facta est grando 8:7 The first angel sounded, and there followed hail and

fire mingled with blood, and they were cast upon the earth: and the third part of trees was burnt up, and all green grass

et ignis mixta in sanguine et missum est in terram et tertia was burnt up.

Biblia Sacra Vulgata

pars terrae conbusta est et tertia pars arborum conbusta est et

8:8 And the second angel sounded, and as it were a great omne faenum uiride conbustum est 8:8 et secundus ange-

mountain burning with fire was cast into the sea: and the third part of the sea became blood;

lus tuba cecinit et tamquam mons magnus igne ardens mis-

8:9 And the third part of the creatures which were in the sus est in mare et facta est tertia pars maris sanguis 8:9 et

sea, and had life, died; and the third part of the ships were destroyed.

mortua est tertia pars creaturae quae habent animas et tertia

8:10 And the third angel sounded, and there fell a great pars nauium interiit 8:10 et tertius angelus tuba cecinit et

star from heaven, burning as it were a lamp, and it fell upon the third part of the rivers, and upon the fountains of

waters; cecidit de caelo stella magna ardens tamquam facula et ceci-8:11 And the name of the star is called Wormwood: and dit in tertiam partem fluminum et in fontes aquarum 8:11 et

the third part of the waters became wormwood; and many men died of the waters, because they were made bitter.

nomen stellae dicitur Absinthius et facta est tertia pars aquarum in absinthium et multi hominum mortui sunt de aquis

8:12 And the fourth angel sounded, and the third part of quia amarae factae sunt 8:12 et quartus angelus tuba ceci-

the sun was smitten, and the third part of the moon, and the third part of the stars; so as the third part of them was darkened, and the day shone not for a third part of it, and the night likewise.

nit et percussa est tertia pars solis et tertia pars lunae et tertia pars stellarum ut obscuraretur tertia pars eorum et diei non

8:13 And I beheld, and heard an angel flying through the luceret pars tertia et nox similiter 8:13 et uidi et audiui uo-

midst of heaven, saying with a loud voice, Woe, woe, woe, to the inhabiters of the earth by reason of the other voices

of the trumpet of the three angels, which are yet to sound! cem unius aquilae uolantis per medium caelum dicentis uoce magna uae uae uae habitantibus in terra de ceteris uocibus tubae trium angelorum qui erant tuba canituri

9:1 And the fifth angel sounded, and I saw a star fall from 9:1 et quintus angelus tuba cecinit et uidi stellam de caelo

heaven unto the earth: and to him was given the key of the bottomless pit.

cecidisse in terram et data est illi clauis putei abyssi 9:2

9:2 And he opened the bottomless pit; and there arose a smoke out of the pit, as the smoke of a great furnace; and the sun and the air were darkened by reason of the smoke of the pit.

et aperuit puteum abyssi et ascendit fumus putei sicut fumus fornacis magnae et obscuratus est sol et aer de fumo pu-

9:3 And there came out of the smoke locusts upon the tei 9:3 et de fumo exierunt lucustae in terram et data est il-

earth: and unto them was given power, as the scorpions of the earth have power.

lis potestas sicut habent potestatem scorpiones terrae 9:4 et

9:4 And it was commanded them that they should not hurt the grass of the earth, neither any green thing, neither any tree; but only those men which have not the seal of God in their foreheads.

praeceptum est illis ne laederent faenum terrae neque omne uiride neque omnem arborem nisi tantum homines qui non

9:5 And to them it was given that they should not kill them, habent signum Dei in frontibus 9:5 et datum est illis ne oc-

but that they should be tormented five months: and their torment was as the torment of a scorpion, when he striketh

a man. ciderent eos sed ut cruciarentur mensibus quinque et crucia-9:6 And in those days shall men seek death, and shall not tus eorum ut cruciatus scorpii cum percutit hominem 9:6

find it; and shall desire to die, and death shall flee from them.

et in diebus illis quaerent homines mortem et non inuenient

9:7 And the shapes of the locusts were like unto horses eam et desiderabunt mori et fugiet mors ab ipsis 9:7 et si-

prepared unto battle; and on their heads were as it were crowns like gold, and their faces were as the faces of men.

militudines lucustarum similes equis paratis in proelium et super capita earum tamquam coronae similes auro et facies

9:8 And they had hair as the hair of women, and their teeth earum sicut facies hominum 9:8 et habebant capillos sicut

were as the teeth of lions.

Apocalypsis Ionannis

capillos mulierum et dentes earum sicut leonum erant 9:9 9:9 And they had breastplates, as it were breastplates of

iron; and the sound of their wings was as the sound of chariots of many horses running to battle.

et habebant loricas sicut loricas ferreas et uox alarum earum sicut uox curruum equorum multorum currentium in bellum 9:10 et habebant caudas similes scorpionum et aculei in cau- 9:10 And they had tails like unto scorpions, and there were

stings in their tails: and their power was to hurt men five months.

dis earum potestas earum nocere hominibus mensibus quinque 9:11 et habebant super se regem angelum abyssi cui no- 9:11 And they had a king over them, which is the angel

of the bottomless pit, whose name in the Hebrew tongue is Abaddon, but in the Greek tongue hath his name Apol-

men hebraice Abaddon graece autem Apollyon et latine ha- lyon. bet nomen Exterminans 9:12 uae unum abiit ecce ueniunt 9:12 One woe is past; and, behold, there come two woes

more hereafter.

adhuc duo uae post haec 9:13 et sextus angelus tuba cecinit 9:13 And the sixth angel sounded, and I heard a voice from

the four horns of the golden altar which is before God,

et audiui uocem unum ex cornibus altaris aurei quod est ante oculos Dei 9:14 dicentem sexto angelo qui habebat tubam 9:14 Saying to the sixth angel which had the trumpet,

Loose the four angels which are bound in the great river Euphrates.

solue quattuor angelos qui alligati sunt in flumine magno Eufrate 9:15 et soluti sunt quattuor angeli qui parati erant 9:15 And the four angels were loosed, which were pre-

pared for an hour, and a day, and a month, and a year, for to slay the third part of men.

in horam et diem et mensem et annum ut occiderent tertiam partem hominum 9:16 et numerus equestris exercitus uicies 9:16 And the number of the army of the horsemen were

two hundred thousand thousand: and I heard the number of them.

milies dena milia audiui numerum eorum 9:17 et ita uidi

9:17 And thus I saw the horses in the vision, and them that sat on them, having breastplates of fire, and of jacinth, and brimstone: and the heads of the horses were as the heads of lions; and out of their mouths issued fire and smoke and brimstone.

equos in uisione et qui sedebant super eos habentes loricas igneas et hyacinthinas et sulphureas et capita equorum erant tamquam capita leonum et de ore ipsorum procedit ignis et fumus et sulphur 9:18 ab his tribus plagis occisa est tertia 9:18 By these three was the third part of men killed, by the

fire, and by the smoke, and by the brimstone, which issued out of their mouths.

pars hominum de igne et fumo et sulphure qui procedebat ex

ore ipsorum 9:19 potestas enim equorum in ore eorum est 9:19 For their power is in their mouth, and in their tails: for their tails were like unto serpents, and had heads, and with them they do hurt.

et in caudis eorum nam caudae illorum similes serpentibus habentes capita et in his nocent 9:20 et ceteri homines qui 9:20 And the rest of the men which were not killed by

these plagues yet repented not of the works of their hands, that they should not worship devils, and idols of gold, and silver, and brass, and stone, and of wood: which neither can see, nor hear, nor walk:

non sunt occisi in his plagis neque paenitentiam egerunt de operibus manuum suarum ut non adorarent daemonia et simulacra aurea et argentea et aerea et lapidea et lignea quae neque uidere possunt neque audire neque ambulare 9:21 et 9:21 Neither repented they of their murders, nor of their

sorceries, nor of their fornication, nor of their thefts.

non egerunt paenitentiam ab homicidiis suis neque a ueneficiis suis neque a fornicatione sua neque a furtis suis

10:1 et uidi alium angelum fortem descendentem de caelo 10:1 And I saw another mighty angel come down from

heaven, clothed with a cloud: and a rainbow was upon his head, and his face was as it were the sun, and his feet as

amictum nube et iris in capite eius et facies eius erat ut sol pillars of fire: et pedes eius tamquam columna ignis 10:2 et habebat in 10:2 And he had in his hand a little book open: and he set

his right foot upon the sea, and his left foot on the earth,

Biblia Sacra Vulgata

manu sua libellum apertum et posuit pedem suum dextrum

10:3 And cried with a loud voice, as when a lion roareth: supra mare sinistrum autem super terram 10:3 et clamauit

and when he had cried, seven thunders uttered their voices.

uoce magna quemadmodum cum leo rugit et cum clamasset

10:4 And when the seven thunders had uttered their voices, locuta sunt septem tonitrua uoces suas 10:4 et cum locuta I was about to write: and I heard a voice from heaven saying unto me, Seal up those things which the seven thunders

uttered, and write them not. fuissent septem tonitrua scripturus eram et audiui uocem de

caelo dicentem signa quae locuta sunt septem tonitrua et noli

10:5 And the angel which I saw stand upon the sea and ea scribere 10:5 et angelum quem uidi stantem supra mare

upon the earth lifted up his hand to heaven,

10:6 And sware by him that liveth for ever and ever, who et supra terram leuauit manum suam ad caelum 10:6 et iu-

created heaven, and the things that therein are, and the earth, and the things that therein are, and the sea, and the things which are therein, that there should be time no longer:

rauit per uiuentem in saecula saeculorum qui creauit caelum et ea quae in illo sunt et terram et ea quae in ea sunt et mare

10:7 But in the days of the voice of the seventh an- et quae in eo sunt quia tempus amplius non erit 10:7 sed

gel, when he shall begin to sound, the mystery of God should be finished, as he hath declared to his servants the

prophets. in diebus uocis septimi angeli cum coeperit tuba canere et

consummabitur mysterium Dei sicut euangelizauit per ser-

10:8 And the voice which I heard from heaven spake unto uos suos prophetas 10:8 et uox quam audiui de caelo iterum

me again, and said, Go and take the little book which is open in the hand of the angel which standeth upon the sea

and upon the earth. loquentem mecum et dicentem uade accipe librum apertum 10:9 And I went unto the angel, and said unto him, Give de manu angeli stantis supra mare et supra terram 10:9 et

me the little book. And he said unto me, Take it, and eat it up; and it shall make thy belly bitter, but it shall be in thy

mouth sweet as honey. abii ad angelum dicens ei ut daret mihi librum et dicit mihi

accipe et deuora illum et faciet amaricare uentrem tuum sed

10:10 And I took the little book out of the angel’s hand, in ore tuo erit dulce tamquam mel 10:10 et accepi librum de

and ate it up; and it was in my mouth sweet as honey: and as soon as I had eaten it, my belly was bitter.

manu angeli et deuoraui eum et erat in ore meo tamquam mel dulce et cum deuorassem eum amaricatus est uenter meus

10:11 And he said unto me, Thou must prophesy again 10:11 et dicunt mihi oportet te iterum prophetare populis et

before many peoples, and nations, and tongues, and kings.

gentibus et linguis et regibus multis

11:1 And there was given me a reed like unto a rod: and 11:1 et datus est mihi calamus similis uirgae dicens surge

the angel stood, saying, Rise, and measure the temple of God, and the altar, and them that worship therein.

et metire templum Dei et altare et adorantes in eo 11:2 at-

11:2 But the court which is without the temple leave out, and measure it not; for it is given unto the Gentiles: and the holy city shall they tread under foot forty and two months.

rium autem quod est foris templum eice foras et ne metieris eum quoniam datum est gentibus et ciuitatem sanctam calca-

11:3 And I will give power unto my two witnesses, and bunt mensibus quadraginta duobus 11:3 et dabo duobus te-

they shall prophesy a thousand two hundred and threescore days, clothed in sackcloth.

stibus meis et prophetabunt diebus mille ducentis sexaginta

11:4 These are the two olive trees, and the two candle- amicti saccos 11:4 hii sunt duo oliuae et duo candelabra in

sticks standing before the God of the earth.

11:5 And if any man will hurt them, fire proceedeth out of conspectu Domini terrae stantes 11:5 et si quis eos uoluerit

their mouth, and devoureth their enemies: and if any man will hurt them, he must in this manner be killed.

nocere ignis exiet de ore illorum et deuorabit inimicos eorum

11:6 These have power to shut heaven, that it rain not in et si quis uoluerit eos laedere sic oportet eum occidi 11:6 hii

the days of their prophecy: and have power over waters to turn them to blood, and to smite the earth with all plagues, as often as they will.

Apocalypsis Ionannis

habent potestatem cludendi caelum ne pluat diebus prophetiae ipsorum et potestatem habent super aquas conuertendi eas in sanguinem et percutere terram omni plaga quotienscumque uoluerint 11:7 et cum finierint testimonium suum 11:7 And when they shall have finished their testimony,

the beast that ascendeth out of the bottomless pit shall make war against them, and shall overcome them, and kill

bestia quae ascendit de abysso faciet aduersus illos bellum et them. uincet eos et occidet illos 11:8 et corpora eorum in plateis 11:8 And their dead bodies shall lie in the street of the

great city, which spiritually is called Sodom and Egypt, where also our Lord was crucified.

ciuitatis magnae quae uocatur spiritaliter Sodoma et Aegyptus ubi et Dominus eorum crucifixus est 11:9 et uidebunt de 11:9 And they of the people and kindreds and tongues and

nations shall see their dead bodies three days and an half, and shall not suffer their dead bodies to be put in graves.

populis et tribubus et linguis et gentibus corpora eorum per tres dies et dimidium et corpora eorum non sinunt poni in monumentis 11:10 et inhabitantes terram gaudebunt super 11:10 And they that dwell upon the earth shall rejoice over

them, and make merry, and shall send gifts one to another; because these two prophets tormented them that dwelt on

illis et iucundabuntur et munera mittent inuicem quoniam the earth. hii duo prophetae cruciauerunt eos qui inhabitant super terram 11:11 et post dies tres et dimidium spiritus uitae a Deo 11:11 And after three days and an half the spirit of life

from God entered into them, and they stood upon their feet; and great fear fell upon them which saw them.

intrauit in eos et steterunt super pedes suos et timor magnus cecidit super eos qui uiderunt eos 11:12 et audierunt uocem 11:12 And they heard a great voice from heaven saying

unto them, Come up hither. And they ascended up to heaven in a cloud; and their enemies beheld them.

magnam de caelo dicentem illis ascendite huc et ascenderunt in caelum in nube et uiderunt illos inimici eorum 11:13 et in 11:13 And the same hour was there a great earthquake,

and the tenth part of the city fell, and in the earthquake were slain of men seven thousand: and the remnant were

illa hora factus est terraemotus magnus et decima pars ciui- affrighted, and gave glory to the God of heaven. tatis cecidit et occisi sunt in terraemotu nomina hominum septem milia et reliqui in timore sunt missi et dederunt gloriam Deo caeli 11:14 uae secundum abiit ecce uae tertium ueniet 11:14 The second woe is past; and, behold, the third woe

cometh quickly.

cito 11:15 et septimus angelus tuba cecinit et factae sunt uo- 11:15 And the seventh angel sounded; and there were

great voices in heaven, saying, The kingdoms of this world are become the kingdoms of our Lord, and of his Christ;

ces magnae in caelo dicentes factum est regnum huius mundi and he shall reign for ever and ever. Domini nostri et Christi eius et regnabit in saecula saeculorum 11:16 et uiginti quattuor seniores qui in conspectu Dei 11:16 And the four and twenty elders, which sat before

God on their seats, fell upon their faces, and worshipped God,

sedent in sedibus suis ceciderunt in facies suas et adorauerunt Deum 11:17 dicentes gratias agimus tibi Domine Deus 11:17 Saying, We give thee thanks, O LORD God

Almighty, which art, and wast, and art to come; because thou hast taken to thee thy great power, and hast reigned.

omnipotens qui es et qui eras quia accepisti uirtutem tuam magnam et regnasti 11:18 et iratae sunt gentes et aduenit 11:18 And the nations were angry, and thy wrath is come,

and the time of the dead, that they should be judged, and that thou shouldest give reward unto thy servants the prophets, and to the saints, and them that fear thy name, small and great; and shouldest destroy them which destroy the earth.

ira tua et tempus mortuorum iudicari et reddere mercedem seruis tuis prophetis et sanctis et timentibus nomen tuum pusillis et magnis et exterminandi eos qui corruperunt terram

Biblia Sacra Vulgata

11:19 And the temple of God was opened in heaven, and 11:19 et apertum est templum Dei in caelo et uisa est arca

there was seen in his temple the ark of his testament: and there were lightnings, and voices, and thunderings, and an

earthquake, and great hail. testamenti eius in templo eius et facta sunt fulgora et uoces et terraemotus et grando magna

12:1 And there appeared a great wonder in heaven; a 12:1 et signum magnum paruit in caelo mulier amicta sole

woman clothed with the sun, and the moon under her feet, and upon her head a crown of twelve stars:

et luna sub pedibus eius et in capite eius corona stellarum

12:2 And she being with child cried, travailing in birth, duodecim 12:2 et in utero habens et clamat parturiens et and pained to be delivered.

12:3 And there appeared another wonder in heaven; and cruciatur ut pariat 12:3 et uisum est aliud signum in caelo

behold a great red dragon, having seven heads and ten horns, and seven crowns upon his heads.

et ecce draco magnus rufus habens capita septem et cornua

12:4 And his tail drew the third part of the stars of heaven, decem et in capitibus suis septem diademata 12:4 et cauda and did cast them to the earth: and the dragon stood before the woman which was ready to be delivered, for to devour

her child as soon as it was born. eius trahebat tertiam partem stellarum caeli et misit eas in

terram et draco stetit ante mulierem quae erat paritura ut cum

12:5 And she brought forth a man child, who was to rule peperisset filium eius deuoraret 12:5 et peperit filium mas-

all nations with a rod of iron: and her child was caught up unto God, and to his throne.

culum qui recturus erit omnes gentes in uirga ferrea et raptus

12:6 And the woman fled into the wilderness, where she est filius eius ad Deum et ad thronum eius 12:6 et mulier fu-

hath a place prepared of God, that they should feed her there a thousand two hundred and threescore days.

git in solitudinem ubi habet locum paratum a Deo ut ibi pa-

12:7 And there was war in heaven: Michael and his angels scant illam diebus mille ducentis sexaginta 12:7 et factum

fought against the dragon; and the dragon fought and his angels,

est proelium in caelo Michahel et angeli eius proeliabantur

12:8 And prevailed not; neither was their place found any cum dracone et draco pugnabat et angeli eius 12:8 et non more in heaven.

ualuerunt neque locus inuentus est eorum amplius in caelo

12:9 And the great dragon was cast out, that old serpent, 12:9 et proiectus est draco ille magnus serpens antiquus qui

called the Devil, and Satan, which deceiveth the whole world: he was cast out into the earth, and his angels were

cast out with him. uocatur Diabolus et Satanas qui seducit uniuersum orbem

proiectus est in terram et angeli eius cum illo missi sunt

12:10 And I heard a loud voice saying in heaven, Now is 12:10 et audiui uocem magnam in caelo dicentem nunc facta

come salvation, and strength, and the kingdom of our God, and the power of his Christ: for the accuser of our brethren is cast down, which accused them before our God day and night.

est salus et uirtus et regnum Dei nostri et potestas Christi eius quia proiectus est accusator fratrum nostrorum qui ac-

12:11 And they overcame him by the blood of the Lamb, cusabat illos ante conspectum Dei nostri die ac nocte 12:11

and by the word of their testimony; and they loved not their lives unto the death.

et ipsi uicerunt illum propter sanguinem agni et propter uerbum testimonii sui et non dilexerunt animam suam usque ad

12:12 Therefore rejoice, ye heavens, and ye that dwell in mortem 12:12 propterea laetamini caeli et qui habitatis in

them. Woe to the inhabiters of the earth and of the sea! for the devil is come down unto you, having great wrath,

because he knoweth that he hath but a short time. eis uae terrae et mari quia descendit diabolus ad uos habens 12:13 And when the dragon saw that he was cast unto the iram magnam sciens quod modicum tempus habet 12:13 et

earth, he persecuted the woman which brought forth the man child.

postquam uidit draco quod proiectus est in terram persecu-

12:14 And to the woman were given two wings of a great tus est mulierem quae peperit masculum 12:14 et datae sunt

eagle, that she might fly into the wilderness, into her place, where she is nourished for a time, and times, and half a time, from the face of the serpent.

Apocalypsis Ionannis

mulieri duae alae aquilae magnae ut uolaret in desertum in locum suum ubi alitur per tempus et tempora et dimidium temporis a facie serpentis 12:15 et misit serpens ex ore suo 12:15 And the serpent cast out of his mouth water as a

flood after the woman, that he might cause her to be carried away of the flood.

post mulierem aquam tamquam flumen ut eam faceret trahi a flumine 12:16 et adiuuit terra mulierem et aperuit terra os 12:16 And the earth helped the woman, and the earth

opened her mouth, and swallowed up the flood which the dragon cast out of his mouth.

suum et absorbuit flumen quod misit draco de ore suo 12:17

12:17 And the dragon was wroth with the woman, and went to make war with the remnant of her seed, which keep the commandments of God, and have the testimony of Jesus Christ.

et iratus est draco in mulierem et abiit facere proelium cum reliquis de semine eius qui custodiunt mandata Dei et habent testimonium Iesu 12:18 et steti super harenam maris 12:18 And I stood upon the sand of the sea,

13:1 et uidi de mare bestiam ascendentem habentem capita 13:1 and saw a beast rise up out of the sea, having seven

heads and ten horns, and upon his horns ten crowns, and upon his heads the name of blasphemy.

septem et cornua decem et super cornua eius decem diademata et super capita eius nomina blasphemiae 13:2 et bes- 13:2 And the beast which I saw was like unto a leopard,

and his feet were as the feet of a bear, and his mouth as the mouth of a lion: and the dragon gave him his power, and

tiam quam uidi similis erat pardo et pedes eius sicut ursi et os his seat, and great authority. eius sicut os leonis et dedit illi draco uirtutem suam et potestatem magnam 13:3 et unum de capitibus suis quasi occi- 13:3 And I saw one of his heads as it were wounded to

death; and his deadly wound was healed: and all the world wondered after the beast.

sum in mortem et plaga mortis eius curata est et admirata est uniuersa terra post bestiam 13:4 et adorauerunt draconem 13:4 And they worshipped the dragon which gave power

unto the beast: and they worshipped the beast, saying, Who is like unto the beast? who is able to make war with

quia dedit potestatem bestiae et adorauerunt bestiam dicen- him? tes quis similis bestiae et quis poterit pugnare cum ea 13:5 13:5 And there was given unto him a mouth speaking great

things and blasphemies; and power was given unto him to continue forty and two months.

et datum est ei os loquens magna et blasphemiae et data est illi potestas facere menses quadraginta duo 13:6 et aperuit 13:6 And he opened his mouth in blasphemy against God,

to blaspheme his name, and his tabernacle, and them that dwell in heaven.

os suum in blasphemias ad Deum blasphemare nomen eius et tabernaculum eius et eos qui in caelo habitant 13:7 et da- 13:7 And it was given unto him to make war with the

saints, and to overcome them: and power was given him over all kindreds, and tongues, and nations.

tum est illi bellum facere cum sanctis et uincere illos et data est ei potestas in omnem tribum et populum et linguam et gentem 13:8 et adorabunt eum omnes qui inhabitant terram 13:8 And all that dwell upon the earth shall worship him,

whose names are not written in the book of life of the Lamb slain from the foundation of the world.

quorum non sunt scripta nomina in libro uitae agni qui occisus est ab origine mundi 13:9 si quis habet aurem audiat 13:9 If any man have an ear, let him hear. 13:10 qui in captiuitatem in captiuitatem uadit qui in gladio 13:10 He that leadeth into captivity shall go into captiv-

ity: he that killeth with the sword must be killed with the sword. Here is the patience and the faith of the saints.

occiderit oportet eum gladio occidi hic est patientia et fides sanctorum 13:11 et uidi aliam bestiam ascendentem de 13:11 And I beheld another beast coming up out of the

earth; and he had two horns like a lamb, and he spake as a dragon.

terra et habebat cornua duo similia agni et loquebatur sicut draco 13:12 et potestatem prioris bestiae omnem faciebat in 13:12 And he exerciseth all the power of the first beast

before him, and causeth the earth and them which dwell therein to worship the first beast, whose deadly wound was healed.

Biblia Sacra Vulgata

conspectu eius et facit terram et inhabitantes in eam adorare

13:13 And he doeth great wonders, so that he maketh fire bestiam primam cuius curata est plaga mortis 13:13 et fecit

come down from heaven on the earth in the sight of men,

signa magna ut etiam ignem faceret de caelo descendere in

13:14 And deceiveth them that dwell on the earth by the terram in conspectu hominum 13:14 et seducit habitantes

means of those miracles which he had power to do in the sight of the beast; saying to them that dwell on the earth, that they should make an image to the beast, which had the wound by a sword, and did live.

terram propter signa quae data sunt illi facere in conspectu bestiae dicens habitantibus in terra ut faciant imaginem be-

13:15 And he had power to give life unto the image of the stiae quae habet plagam gladii et uixit 13:15 et datum est

beast, that the image of the beast should both speak, and cause that as many as would not worship the image of the

beast should be killed. illi ut daret spiritum imagini bestiae ut et loquatur imago be-

stiae et faciat quicumque non adorauerint imaginem bestiae

13:16 And he causeth all, both small and great, rich and occidantur 13:16 et faciet omnes pusillos et magnos et diui-

poor, free and bond, to receive a mark in their right hand, or in their foreheads:

tes et pauperes et liberos et seruos habere caracter in dextera

13:17 And that no man might buy or sell, save he that had manu aut in frontibus suis 13:17 et ne quis possit emere aut

the mark, or the name of the beast, or the number of his name.

uendere nisi qui habet caracter nomen bestiae aut numerum

13:18 Here is wisdom. Let him that hath understanding nominis eius 13:18 hic sapientia est qui habet intellectum

count the number of the beast: for it is the number of a man; and his number is Six hundred threescore and six.

conputet numerum bestiae numerus enim hominis est et numerus eius est sescenti sexaginta sex

14:1 And I looked, and, lo, a Lamb stood on the mount 14:1 et uidi et ecce agnus stabat supra montem Sion et cum

Sion, and with him an hundred forty and four thousand, having his Father’s name written in their foreheads.

illo centum quadraginta quattuor milia habentes nomen eius

14:2 And I heard a voice from heaven, as the voice of et nomen Patris eius scriptum in frontibus suis 14:2 et au-

many waters, and as the voice of a great thunder: and I heard the voice of harpers harping with their harps:

diui uocem de caelo tamquam uocem aquarum multarum et tamquam uocem tonitrui magni et uocem quam audiui sicut

14:3 And they sung as it were a new song before the citharoedorum citharizantium in citharis suis 14:3 et canta-

throne, and before the four beasts, and the elders: and no man could learn that song but the hundred and forty and

four thousand, which were redeemed from the earth. bant quasi canticum nouum ante sedem et ante quattuor animalia et seniores et nemo poterat discere canticum nisi illa centum quadraginta quattuor milia qui empti sunt de terra

14:4 These are they which were not defiled with women; 14:4 hii sunt qui cum mulieribus non sunt coinquinati uir-

for they are virgins. These are they which follow the Lamb whithersoever he goeth. These were redeemed from among men, being the firstfruits unto God and to the Lamb.

gines enim sunt hii qui sequuntur agnum quocumque abierit

14:5 And in their mouth was found no guile: for they are hii empti sunt ex hominibus primitiae Deo et agno 14:5 et in

without fault before the throne of God.

ore ipsorum non est inuentum mendacium sine macula sunt

14:6 And I saw another angel fly in the midst of heaven, 14:6 et uidi alterum angelum uolantem per medium caelum

having the everlasting gospel to preach unto them that dwell on the earth, and to every nation, and kindred, and

tongue, and people, habentem euangelium aeternum ut euangelizaret sedentibus

super terram et super omnem gentem et tribum et linguam

14:7 Saying with a loud voice, Fear God, and give glory et populum 14:7 dicens magna uoce timete Deum et date to him; for the hour of his judgment is come: and worship him that made heaven, and earth, and the sea, and the fountains of waters.

Apocalypsis Ionannis

illi honorem quia uenit hora iudicii eius et adorate eum qui

fecit caelum et terram et mare et fontes aquarum 14:8 et 14:8 And there followed another angel, saying, Babylon is fallen, is fallen, that great city, because she made all nations drink of the wine of the wrath of her fornication.

alius angelus secutus est dicens cecidit cecidit Babylon illa magna quae a uino irae fornicationis suae potionauit omnes gentes 14:9 et alius angelus tertius secutus est illos dicens 14:9 And the third angel followed them, saying with a loud

voice, If any man worship the beast and his image, and receive his mark in his forehead, or in his hand,

uoce magna si quis adorauerit bestiam et imaginem eius et acceperit caracterem in fronte sua aut in manu sua 14:10 et 14:10 The same shall drink of the wine of the wrath of

God, which is poured out without mixture into the cup of his indignation; and he shall be tormented with fire and brimstone in the presence of the holy angels, and in the presence of the Lamb:

hic bibet de uino irae Dei qui mixtus est mero in calice irae ipsius et cruciabitur igne et sulphure in conspectu angelorum sanctorum et ante conspectum agni 14:11 et fumus tor- 14:11 And the smoke of their torment ascendeth up for

ever and ever: and they have no rest day nor night, who worship the beast and his image, and whosoever receiveth

mentorum eorum in saecula saeculorum ascendit nec habent the mark of his name. requiem die ac nocte qui adorauerunt bestiam et imaginem eius et si quis acceperit caracterem nominis eius 14:12 hic 14:12 Here is the patience of the saints: here are they that

keep the commandments of God, and the faith of Jesus.

patientia sanctorum est qui custodiunt mandata Dei et fidem Iesu 14:13 et audiui uocem de caelo dicentem scribe beati 14:13 And I heard a voice from heaven saying unto me,

Write, Blessed are the dead which die in the Lord from henceforth: Yea, saith the Spirit, that they may rest from

mortui qui in Domino moriuntur amodo iam dicit Spiritus their labours; and their works do follow them. ut requiescant a laboribus suis opera enim illorum sequuntur illos 14:14 et uidi et ecce nubem candidam et supra nubem 14:14 And I looked, and behold a white cloud, and upon

the cloud one sat like unto the Son of man, having on his head a golden crown, and in his hand a sharp sickle.

sedentem similem Filio hominis habentem in capite suo coronam auream et in manu sua falcem acutam 14:15 et alter 14:15 And another angel came out of the temple, crying

with a loud voice to him that sat on the cloud, Thrust in thy sickle, and reap: for the time is come for thee to reap;

angelus exiuit de templo clamans uoce magna ad sedentem for the harvest of the earth is ripe. super nubem mitte falcem tuam et mete quia uenit hora ut metatur quoniam aruit messis terrae 14:16 et misit qui se- 14:16 And he that sat on the cloud thrust in his sickle on

the earth; and the earth was reaped.

debat supra nubem falcem suam in terram et messa est terra 14:17 et alius angelus exiuit de templo quod est in caelo ha- 14:17 And another angel came out of the temple which is

in heaven, he also having a sharp sickle.

bens et ipse falcem acutam 14:18 et alius angelus de altari 14:18 And another angel came out from the altar, which

had power over fire; and cried with a loud cry to him that had the sharp sickle, saying, Thrust in thy sharp sickle, and gather the clusters of the vine of the earth; for her grapes are fully ripe.

qui habet potestatem supra ignem et clamauit uoce magna qui habebat falcem acutam dicens mitte falcem tuam acutam et uindemia botros uineae terrae quoniam maturae sunt uuae

eius 14:19 et misit angelus falcem suam in terram et uin- 14:19 And the angel thrust in his sickle into the earth, and gathered the vine of the earth, and cast it into the great winepress of the wrath of God.

demiauit uineam terrae et misit in lacum irae Dei magnum 14:20 et calcatus est lacus extra ciuitatem et exiuit sanguis de 14:20 And the winepress was trodden without the city, and

blood came out of the winepress, even unto the horse bridles, by the space of a thousand and six hundred furlongs.

lacu usque ad frenos equorum per stadia mille sescenta

Biblia Sacra Vulgata

15:1 And I saw another sign in heaven, great and marvel- 15:1 et uidi aliud signum in caelo magnum et mirabile an-

lous, seven angels having the seven last plagues; for in them is filled up the wrath of God.

gelos septem habentes plagas septem nouissimas quoniam in

15:2 And I saw as it were a sea of glass mingled with fire: illis consummata est ira Dei 15:2 et uidi tamquam mare ui-

and them that had gotten the victory over the beast, and over his image, and over his mark, and over the number of his name, stand on the sea of glass, having the harps of God.

treum mixtum igne et eos qui uicerunt bestiam et imaginem illius et numerum nominis eius stantes supra mare uitreum

15:3 And they sing the song of Moses the servant of God, habentes citharas Dei 15:3 et cantant canticum Mosi serui

and the song of the Lamb, saying, Great and marvellous are thy works, Lord God Almighty; just and true are thy

ways, thou King of saints. Dei et canticum agni dicentes magna et mirabilia opera tua

Domine Deus omnipotens iustae et uerae uiae tuae rex sae-

15:4 Who shall not fear thee, O Lord, and glorify thy culorum 15:4 quis non timebit Domine et magnificabit no-

name? for thou only art holy: for all nations shall come and worship before thee; for thy judgments are made man-

ifest. men tuum quia solus pius quoniam omnes gentes uenient et

adorabunt in conspectu tuo quoniam iudicia tua manifestata

15:5 And after that I looked, and, behold, the temple of sunt 15:5 et post haec uidi et ecce apertum est templum ta-

the tabernacle of the testimony in heaven was opened:

15:6 And the seven angels came out of the temple, having bernaculi testimonii in caelo 15:6 et exierunt septem angeli

the seven plagues, clothed in pure and white linen, and having their breasts girded with golden girdles.

habentes septem plagas de templo uestiti lapide mundo can-

15:7 And one of the four beasts gave unto the seven angels dido et praecincti circa pectora zonis aureis 15:7 et unus ex

seven golden vials full of the wrath of God, who liveth for ever and ever.

quattuor animalibus dedit septem angelis septem fialas aureas plenas iracundiae Dei uiuentis in saecula saeculorum

15:8 And the temple was filled with smoke from the glory 15:8 et impletum est templum fumo a maiestate Dei et de of God, and from his power; and no man was able to enter into the temple, till the seven plagues of the seven angels

were fulfilled. uirtute eius et nemo poterat introire in templum donec consummarentur septem plagae septem angelorum

16:1 And I heard a great voice out of the temple saying to 16:1 et audiui uocem magnam de templo dicentem septem the seven angels, Go your ways, and pour out the vials of the wrath of God upon the earth.

angelis ite et effundite septem fialas irae Dei in terram 16:2

16:2 And the first went, and poured out his vial upon the earth; and there fell a noisome and grievous sore upon the men which had the mark of the beast, and upon them which worshipped his image.

et abiit primus et effudit fialam suam in terram et factum est uulnus saeuum ac pessimum in homines qui habent carac-

16:3 And the second angel poured out his vial upon the terem bestiae et eos qui adorauerunt imaginem eius 16:3

sea; and it became as the blood of a dead man: and every living soul died in the sea.

et secundus effudit fialam suam in mare et factus est sanguis tamquam mortui et omnis anima uiuens mortua est in

16:4 And the third angel poured out his vial upon the rivers mari 16:4 et tertius effudit fialam suam super flumina et suand fountains of waters; and they became blood.

16:5 And I heard the angel of the waters say, Thou art righ- per fontes aquarum et factus est sanguis 16:5 et audiui an-

teous, O Lord, which art, and wast, and shalt be, because thou hast judged thus.

gelum aquarum dicentem iustus es qui es et qui eras sanc-

16:6 For they have shed the blood of saints and prophets, tus quia haec iudicasti 16:6 quia sanguinem sanctorum et

and thou hast given them blood to drink; for they are worthy.

prophetarum fuderunt et sanguinem eis dedisti bibere digni

16:7 And I heard another out of the altar say, Even so, sunt 16:7 et audiui altare dicens etiam Domine Deus om-

Lord God Almighty, true and righteous are thy judgments.

Apocalypsis Ionannis

nipotens uera et iusta iudicia tua 16:8 et quartus effudit fia- 16:8 And the fourth angel poured out his vial upon the sun;

and power was given unto him to scorch men with fire.

lam suam in solem et datum est illi aestu adficere homines et igni 16:9 et aestuauerunt homines aestu magno et blasphe- 16:9 And men were scorched with great heat, and blas-

phemed the name of God, which hath power over these plagues: and they repented not to give him glory.

mauerunt nomen Dei habentis potestatem super has plagas neque egerunt paenitentiam ut darent illi gloriam 16:10 et 16:10 And the fifth angel poured out his vial upon the seat

of the beast; and his kingdom was full of darkness; and they gnawed their tongues for pain,

quintus effudit fialam suam super sedem bestiae et factum est regnum eius tenebrosum et conmanducauerunt linguas suas prae dolore 16:11 et blasphemauerunt Deum caeli prae 16:11 And blasphemed the God of heaven because of their

pains and their sores, and repented not of their deeds.

doloribus et uulneribus suis et non egerunt paenitentiam ex operibus suis 16:12 et sextus effudit fialam suam in flumen 16:12 And the sixth angel poured out his vial upon the

great river Euphrates; and the water thereof was dried up, that the way of the kings of the east might be prepared.

illud magnum Eufraten et siccauit aquam eius ut praepararetur uia regibus ab ortu solis 16:13 et uidi de ore draconis 16:13 And I saw three unclean spirits like frogs come out

of the mouth of the dragon, and out of the mouth of the beast, and out of the mouth of the false prophet.

et de ore bestiae et de ore pseudoprophetae spiritus tres in-

mundos in modum ranarum 16:14 sunt enim spiritus dae- 16:14 For they are the spirits of devils, working miracles, which go forth unto the kings of the earth and of the whole world, to gather them to the battle of that great day of God

moniorum facientes signa et procedunt ad reges totius terrae Almighty. congregare illos in proelium ad diem magnum Dei omnipotentis 16:15 ecce uenio sicut fur beatus qui uigilat et custo- 16:15 Behold, I come as a thief. Blessed is he that watch-

eth, and keepeth his garments, lest he walk naked, and they see his shame.

dit uestimenta sua ne nudus ambulet et uideant turpitudinem eius 16:16 et congregauit illos in locum qui uocatur hebra- 16:16 And he gathered them together into a place called in

the Hebrew tongue Armageddon.

ice Hermagedon 16:17 et septimus effudit fialam suam in 16:17 And the seventh angel poured out his vial into the

air; and there came a great voice out of the temple of heaven, from the throne, saying, It is done.

aerem et exiuit uox magna de templo a throno dicens factum est 16:18 et facta sunt fulgora et uoces et tonitrua et terrae- 16:18 And there were voices, and thunders, and light-

nings; and there was a great earthquake, such as was not since men were upon the earth, so mighty an earthquake,

motus factus est magnus qualis numquam fuit ex quo homi- and so great. nes fuerunt super terram talis terraemotus sic magnus 16:19 16:19 And the great city was divided into three parts, and

the cities of the nations fell: and great Babylon came in remembrance before God, to give unto her the cup of the

et facta est ciuitas magna in tres partes et ciuitates gentium wine of the fierceness of his wrath. ceciderunt et Babylon magna uenit in memoriam ante Deum dare ei calicem uini indignationis irae eius 16:20 et om- 16:20 And every island fled away, and the mountains were

not found.

nis insula fugit et montes non sunt inuenti 16:21 et grando 16:21 And there fell upon men a great hail out of heaven,

every stone about the weight of a talent: and men blasphemed God because of the plague of the hail; for the

magna sicut talentum descendit de caelo in homines et blas- plague thereof was exceeding great. phemauerunt homines Deum propter plagam grandinis quoniam magna facta est uehementer

17:1 et uenit unus de septem angelis qui habebant septem 17:1 And there came one of the seven angels which had

the seven vials, and talked with me, saying unto me, Come hither; I will shew unto thee the judgment of the great

fialas et locutus est mecum dicens ueni ostendam tibi dam- whore that sitteth upon many waters:

Biblia Sacra Vulgata

nationem meretricis magnae quae sedet super aquas multas

17:2 With whom the kings of the earth have committed 17:2 cum qua fornicati sunt reges terrae et inebriati sunt qui

fornication, and the inhabitants of the earth have been made drunk with the wine of her fornication.

inhabitant terram de uino prostitutionis eius 17:3 et abstulit

17:3 So he carried me away in the spirit into the wilderness: and I saw a woman sit upon a scarlet coloured beast, full of names of blasphemy, having seven heads and ten horns.

me in desertum in spiritu et uidi mulierem sedentem super bestiam coccineam plenam nominibus blasphemiae haben-

17:4 And the woman was arrayed in purple and scarlet tem capita septem et cornua decem 17:4 et mulier erat cir-

colour, and decked with gold and precious stones and pearls, having a golden cup in her hand full of abomina-

tions and filthiness of her fornication: cumdata purpura et coccino et inaurata auro et lapide pre-

tioso et margaritis habens poculum aureum in manu sua ple-

17:5 And upon her forehead was a name written, MYS- num abominationum et inmunditia fornicationis eius 17:5

TERY, BABYLON THE GREAT, THE MOTHER OF HARLOTS AND ABOMINATIONS OF THE EARTH.

et in fronte eius nomen scriptum mysterium Babylon magna

17:6 And I saw the woman drunken with the blood of the mater fornicationum et abominationum terrae 17:6 et uidi

saints, and with the blood of the martyrs of Jesus: and when I saw her, I wondered with great admiration.

mulierem ebriam de sanguine sanctorum et de sanguine martyrum Iesu et miratus sum cum uidissem illam admiratione

17:7 And the angel said unto me, Wherefore didst thou magna 17:7 et dixit mihi angelus quare miraris ego tibi di-

marvel? I will tell thee the mystery of the woman, and of the beast that carrieth her, which hath the seven heads and

ten horns. cam sacramentum mulieris et bestiae quae portat eam quae

17:8 The beast that thou sawest was, and is not; and shall habet capita septem et decem cornua 17:8 bestiam quam ascend out of the bottomless pit, and go into perdition: and they that dwell on the earth shall wonder, whose names were not written in the book of life from the foundation of the world, when they behold the beast that was, and is not, and yet is.

uidisti fuit et non est et ascensura est de abysso et in interitum ibit et mirabuntur inhabitantes terram quorum non sunt scripta nomina in libro uitae a constitutione mundi uiden-

17:9 And here is the mind which hath wisdom. The seven tes bestiam quia erat et non est 17:9 et hic est sensus qui

heads are seven mountains, on which the woman sitteth.

habet sapientiam septem capita septem montes sunt super

17:10 And there are seven kings: five are fallen, and one quos mulier sedet et reges septem sunt 17:10 quinque ce-

is, and the other is not yet come; and when he cometh, he must continue a short space.

ciderunt unus est alius nondum uenit et cum uenerit oportet

17:11 And the beast that was, and is not, even he is the illum breue tempus manere 17:11 et bestia quae erat et non

eighth, and is of the seven, and goeth into perdition.

est et ipsa octaua est et de septem est et in interitum uadit

17:12 And the ten horns which thou sawest are ten kings, 17:12 et decem cornua quae uidisti decem reges sunt qui reg-

which have received no kingdom as yet; but receive power as kings one hour with the beast.

num nondum acceperunt sed potestatem tamquam reges una

17:13 These have one mind, and shall give their power and hora accipiunt post bestiam 17:13 hii unum consilium ha-

strength unto the beast.

17:14 These shall make war with the Lamb, and the Lamb bent et uirtutem et potestatem suam bestiae tradunt 17:14

shall overcome them: for he is Lord of lords, and King of kings: and they that are with him are called, and chosen,

and faithful. hii cum agno pugnabunt et agnus uincet illos quoniam Do-

minus dominorum est et rex regum et qui cum illo sunt uo-

17:15 And he saith unto me, The waters which thou cati et electi et fideles 17:15 et dixit mihi aquas quas uidisti

sawest, where the whore sitteth, are peoples, and multitudes, and nations, and tongues.

ubi meretrix sedet populi sunt et gentes et linguae 17:16 et

17:16 And the ten horns which thou sawest upon the beast, these shall hate the whore, and shall make her desolate and naked, and shall eat her flesh, and burn her with fire.

Apocalypsis Ionannis

decem cornua quae uidisti et bestiam hii odient fornicariam et desolatam facient illam et nudam et carnes eius manduca-

bunt et ipsam igni concremabunt 17:17 Deus enim dedit in 17:17 For God hath put in their hearts to fulfil his will, and to agree, and give their kingdom unto the beast, until the words of God shall be fulfilled.

corda eorum ut faciant quod illi placitum est ut dent regnum

suum bestiae donec consummentur uerba Dei 17:18 et mu- 17:18 And the woman which thou sawest is that great city,

which reigneth over the kings of the earth.

lier quam uidisti est ciuitas magna quae habet regnum super reges terrae

18:1 et post haec uidi alium angelum descendentem de 18:1 And after these things I saw another angel come

down from heaven, having great power; and the earth was lightened with his glory.

caelo habentem potestatem magnam et terra inluminata est a gloria eius 18:2 et exclamauit in forti uoce dicens cecidit 18:2 And he cried mightily with a strong voice, saying,

Babylon the great is fallen, is fallen, and is become the habitation of devils, and the hold of every foul spirit, and

cecidit Babylon magna et facta est habitatio daemoniorum et a cage of every unclean and hateful bird. custodia omnis spiritus inmundi et custodia omnis uolucris inmundae 18:3 quia de ira fornicationis eius biberunt om- 18:3 For all nations have drunk of the wine of the wrath of

her fornication, and the kings of the earth have committed fornication with her, and the merchants of the earth are

nes gentes et reges terrae cum illa fornicati sunt et mercato- waxed rich through the abundance of her delicacies. res terrae de uirtute deliciarum eius diuites facti sunt 18:4 et 18:4 And I heard another voice from heaven, saying,

Come out of her, my people, that ye be not partakers of her sins, and that ye receive not of her plagues.

audiui aliam uocem de caelo dicentem exite de illa populus meus ut ne participes sitis delictorum eius et de plagis eius non accipiatis 18:5 quoniam peruenerunt peccata eius us- 18:5 For her sins have reached unto heaven, and God hath

remembered her iniquities.

que ad caelum et recordatus est Deus iniquitatum eius 18:6 18:6 Reward her even as she rewarded you, and double

unto her double according to her works: in the cup which she hath filled fill to her double.

reddite illi sicut ipsa reddidit et duplicate duplicia secundum opera eius in poculo quo miscuit miscite illi duplum 18:7 18:7 How much she hath glorified herself, and lived de-

liciously, so much torment and sorrow give her: for she saith in her heart, I sit a queen, and am no widow, and

quantum glorificauit se et in deliciis fuit tantum date illi tor- shall see no sorrow. mentum et luctum quia in corde suo dicit sedeo regina et

uidua non sum et luctum non uidebo 18:8 ideo in una die 18:8 Therefore shall her plagues come in one day, death, and mourning, and famine; and she shall be utterly burned with fire: for strong is the Lord God who judgeth her.

uenient plagae eius mors et luctus et fames et igni conburetur quia fortis est Deus qui iudicauit illam 18:9 et flebunt 18:9 And the kings of the earth, who have committed for-

nication and lived deliciously with her, shall bewail her, and lament for her, when they shall see the smoke of her

et plangent se super illam reges terrae qui cum illa forni- burning, cati sunt et in deliciis uixerunt cum uiderint fumum incendii eius 18:10 longe stantes propter timorem tormentorum eius 18:10 Standing afar off for the fear of her torment, saying,

Alas, alas that great city Babylon, that mighty city! for in one hour is thy judgment come.

dicentes uae uae ciuitas illa magna Babylon ciuitas illa fortis quoniam una hora uenit iudicium tuum 18:11 et negotiato- 18:11 And the merchants of the earth shall weep and

mourn over her; for no man buyeth their merchandise any more:

res terrae flebunt et lugebunt super illam quoniam merces

eorum nemo emet amplius 18:12 mercem auri et argenti et 18:12 The merchandise of gold, and silver, and precious stones, and of pearls, and fine linen, and purple, and silk, and scarlet, and all thyine wood, and all manner vessels of ivory, and all manner vessels of most precious wood, and of brass, and iron, and marble,

Biblia Sacra Vulgata

lapidis pretiosi et margaritis et byssi et purpurae et serici et cocci et omne lignum thyinum et omnia uasa eboris et omnia uasa de lapide pretioso et aeramento et ferro et marmore

18:13 And cinnamon, and odours, and ointments, and 18:13 et cinnamomum et amomum et odoramentorum et un-

frankincense, and wine, and oil, and fine flour, and wheat, and beasts, and sheep, and horses, and chariots, and slaves,

and souls of men. guenti et turis et uini et olei et similae et tritici et iumen-

torum et ouium et equorum et raedarum et mancipiorum et

18:14 And the fruits that thy soul lusted after are departed animarum hominum 18:14 et poma tua desiderii animae difrom thee, and all things which were dainty and goodly are departed from thee, and thou shalt find them no more

at all. scessit a te et omnia pinguia et clara perierunt a te et amplius 18:15 The merchants of these things, which were made illa iam non inuenient 18:15 mercatores horum qui diuites

rich by her, shall stand afar off for the fear of her torment, weeping and wailing,

facti sunt ab ea longe stabunt propter timorem tormentorum

18:16 And saying, Alas, alas that great city, that was eius flentes ac lugentes 18:16 et dicentes uae uae ciuitas illa

clothed in fine linen, and purple, and scarlet, and decked with gold, and precious stones, and pearls!

magna quae amicta erat byssino et purpura et cocco et deau-

18:17 For in one hour so great riches is come to nought. rata est auro et lapide pretioso et margaritis 18:17 quoniam

And every shipmaster, and all the company in ships, and sailors, and as many as trade by sea, stood afar off,

una hora destitutae sunt tantae diuitiae et omnis gubernator et omnis qui in locum nauigat et nautae et qui maria operan-

18:18 And cried when they saw the smoke of her burning, tur longe steterunt 18:18 et clamauerunt uidentes locum in-

saying, What city is like unto this great city!

18:19 And they cast dust on their heads, and cried, weep- cendii eius dicentes quae similis ciuitati huic magnae 18:19

ing and wailing, saying, Alas, alas that great city, wherein were made rich all that had ships in the sea by reason of

her costliness! for in one hour is she made desolate. et miserunt puluerem super capita sua et clamauerunt flentes et lugentes dicentes uae uae ciuitas magna in qua diuites facti sunt omnes qui habent naues in mari de pretiis eius quo-

18:20 Rejoice over her, thou heaven, and ye holy apostles niam una hora desolata est 18:20 exulta super eam caelum and prophets; for God hath avenged you on her.

et sancti et apostoli et prophetae quoniam iudicauit Deus iu-

18:21 And a mighty angel took up a stone like a great mill- dicium uestrum de illa 18:21 et sustulit unus angelus fortis

stone, and cast it into the sea, saying, Thus with violence shall that great city Babylon be thrown down, and shall be

found no more at all. lapidem quasi molarem magnum et misit in mare dicens hoc

impetu mittetur Babylon magna illa ciuitas et ultra iam non

18:22 And the voice of harpers, and musicians, and of inuenietur 18:22 et uox citharoedorum et musicorum et ti-

pipers, and trumpeters, shall be heard no more at all in thee; and no craftsman, of whatsoever craft he be, shall be found any more in thee; and the sound of a millstone shall be heard no more at all in thee;

bia canentium et tuba non audietur in te amplius et omnis artifex omnis artis non inuenietur in te amplius et uox molae

18:23 And the light of a candle shall shine no more at all non audietur in te amplius 18:23 et lux lucernae non luce-

in thee; and the voice of the bridegroom and of the bride shall be heard no more at all in thee: for thy merchants were the great men of the earth; for by thy sorceries were all nations deceived.

bit tibi amplius et uox sponsi et sponsae non audietur adhuc in te quia mercatores tui erant principes terrae quia in ue-

18:24 And in her was found the blood of prophets, and of neficiis tuis errauerunt omnes gentes 18:24 et in ea sanguis

saints, and of all that were slain upon the earth.

prophetarum et sanctorum inuentus est et omnium qui inter-

Apocalypsis Ionannis

fecti sunt in terra

19:1 post haec audiui quasi uocem magnam turbarum mul- 19:1 And after these things I heard a great voice of much

people in heaven, saying, Alleluia; Salvation, and glory, and honour, and power, unto the Lord our God:

tarum in caelo dicentium alleluia salus et gloria et uirtus Deo nostro est 19:2 quia uera et iusta iudicia sunt eius quia 19:2 For true and righteous are his judgments: for he hath

judged the great whore, which did corrupt the earth with her fornication, and hath avenged the blood of his servants

iudicauit de meretrice magna quae corrupit terram in pros- at her hand. titutione sua et uindicauit sanguinem seruorum suorum de manibus eius 19:3 et iterum dixerunt alleluia et fumus eius 19:3 And again they said, Alleluia And her smoke rose up

for ever and ever.

ascendit in saecula saeculorum 19:4 et ceciderunt senio- 19:4 And the four and twenty elders and the four beasts

fell down and worshipped God that sat on the throne, saying, Amen; Alleluia.

res uiginti quattuor et quattuor animalia et adorauerunt Deum sedentem super thronum dicentes amen alleluia 19:5 et 19:5 And a voice came out of the throne, saying, Praise

our God, all ye his servants, and ye that fear him, both small and great.

uox de throno exiuit dicens laudem dicite Deo nostro omnes serui eius et qui timetis eum pusilli et magni 19:6 et audiui 19:6 And I heard as it were the voice of a great multi-

tude, and as the voice of many waters, and as the voice of mighty thunderings, saying, Alleluia: for the Lord God

quasi uocem turbae magnae et sicut uocem aquarum multa- omnipotent reigneth. rum et sicut uocem tonitruum magnorum dicentium alleluia quoniam regnauit Dominus Deus noster omnipotens 19:7 19:7 Let us be glad and rejoice, and give honour to him:

for the marriage of the Lamb is come, and his wife hath made herself ready.

gaudeamus et exultemus et demus gloriam ei quia uenerunt nuptiae agni et uxor eius praeparauit se 19:8 et datum est 19:8 And to her was granted that she should be arrayed in

fine linen, clean and white: for the fine linen is the righteousness of saints.

illi ut cooperiat se byssinum splendens candidum byssinum enim iustificationes sunt sanctorum 19:9 et dicit mihi scribe 19:9 And he saith unto me, Write, Blessed are they which

are called unto the marriage supper of the Lamb. And he saith unto me, These are the true sayings of God.

beati qui ad cenam nuptiarum agni uocati sunt et dicit mihi haec uerba uera Dei sunt 19:10 et cecidi ante pedes eius ut 19:10 And I fell at his feet to worship him. And he said

unto me, See thou do it not: I am thy fellowservant, and of thy brethren that have the testimony of Jesus: worship

adorarem eum et dicit mihi uide ne feceris conseruus tuus God: for the testimony of Jesus is the spirit of prophecy. sum et fratrum tuorum habentium testimonium Iesu Deum adora testimonium enim Iesu est spiritus prophetiae 19:11 19:11 And I saw heaven opened, and behold a white horse;

and he that sat upon him was called Faithful and True, and in righteousness he doth judge and make war.

et uidi caelum apertum et ecce equus albus et qui sedebat super eum uocabatur Fidelis et Verax uocatur et iustitia iudicat et pugnat 19:12 oculi autem eius sicut flamma ignis 19:12 His eyes were as a flame of fire, and on his head

were many crowns; and he had a name written, that no man knew, but he himself.

et in capite eius diademata multa habens nomen scriptum quod nemo nouit nisi ipse 19:13 et uestitus erat uestem as- 19:13 And he was clothed with a vesture dipped in blood:

and his name is called The Word of God.

persam sanguine et uocatur nomen eius Verbum Dei 19:14 19:14 And the armies which were in heaven followed him

upon white horses, clothed in fine linen, white and clean.

et exercitus qui sunt in caelo sequebantur eum in equis albis uestiti byssinum album mundum 19:15 et de ore ipsius pro- 19:15 And out of his mouth goeth a sharp sword, that with

it he should smite the nations: and he shall rule them with a rod of iron: and he treadeth the winepress of the fierce-

cedit gladius acutus ut in ipso percutiat gentes et ipse reget ness and wrath of Almighty God.

Biblia Sacra Vulgata

eos in uirga ferrea et ipse calcat torcular uini furoris irae Dei

19:16 And he hath on his vesture and on his thigh a name omnipotentis 19:16 et habet in uestimento et in femore suo

written, KING OF KINGS, AND LORD OF LORDS.

19:17 And I saw an angel standing in the sun; and he scriptum rex regum et Dominus dominantium 19:17 et uidi

cried with a loud voice, saying to all the fowls that fly in the midst of heaven, Come and gather yourselves together

unto the supper of the great God; unum angelum stantem in sole et clamauit uoce magna di-

cens omnibus auibus quae uolabant per medium caeli uenite

19:18 That ye may eat the flesh of kings, and the flesh congregamini ad cenam magnam Dei 19:18 ut manducetis

of captains, and the flesh of mighty men, and the flesh of horses, and of them that sit on them, and the flesh of all

men, both free and bond, both small and great. carnes regum et carnes tribunorum et carnes fortium et car-

nes equorum et sedentium in ipsis et carnes omnium libero-

19:19 And I saw the beast, and the kings of the earth, and rum ac seruorum et pusillorum ac magnorum 19:19 et uidi their armies, gathered together to make war against him that sat on the horse, and against his army.

bestiam et reges terrae et exercitus eorum congregatos ad faciendum proelium cum illo qui sedebat in equo et cum ex-

19:20 And the beast was taken, and with him the false ercitu eius 19:20 et adprehensa est bestia et cum illo pseu-

prophet that wrought miracles before him, with which he deceived them that had received the mark of the beast, and them that worshipped his image. These both were cast alive into a lake of fire burning with brimstone.

dopropheta qui fecit signa coram ipso quibus seduxit eos qui acceperunt caracterem bestiae qui et adorant imaginem eius uiui missi sunt hii duo in stagnum ignis ardentis sulphure

19:21 And the remnant were slain with the sword of him 19:21 et ceteri occisi sunt in gladio sedentis super equum qui

that sat upon the horse, which sword proceeded out of his mouth: and all the fowls were filled with their flesh.

procedit de ore ipsius et omnes aues saturatae sunt carnibus eorum

20:1 And I saw an angel come down from heaven, having 20:1 et uidi angelum descendentem de caelo habentem cla-

the key of the bottomless pit and a great chain in his hand.

20:2 And he laid hold on the dragon, that old serpent, uem abyssi et catenam magnam in manu sua 20:2 et adpre-

which is the Devil, and Satan, and bound him a thousand years,

hendit draconem serpentem antiquum qui est diabolus et Sa-

20:3 And cast him into the bottomless pit, and shut him up, tanas et ligauit eum per annos mille 20:3 et misit eum in and set a seal upon him, that he should deceive the nations no more, till the thousand years should be fulfilled: and

after that he must be loosed a little season. abyssum et clusit et signauit super illum ut non seducat am-

plius gentes donec consummentur mille anni post haec opor-

20:4 And I saw thrones, and they sat upon them, and judg- tet illum solui modico tempore 20:4 et uidi sedes et sede-

ment was given unto them: and I saw the souls of them that were beheaded for the witness of Jesus, and for the word of God, and which had not worshipped the beast, neither his image, neither had received his mark upon their foreheads, or in their hands; and they lived and reigned with Christ a thousand years.

runt super eas et iudicium datum est illis et animas decollatorum propter testimonium Iesu et propter uerbum Dei et qui non adorauerunt bestiam neque imaginem eius nec acceperunt caracterem in frontibus aut in manibus suis et uixerunt

20:5 But the rest of the dead lived not again until the thou- et regnauerunt cum Christo mille annis 20:5 ceteri mortuo-

sand years were finished. This is the first resurrection.

rum non uixerunt donec consummentur mille anni haec est

20:6 Blessed and holy is he that hath part in the first resur- resurrectio prima 20:6 beatus et sanctus qui habet partem in

rection: on such the second death hath no power, but they shall be priests of God and of Christ, and shall reign with

him a thousand years. resurrectione prima in his secunda mors non habet potesta-

Apocalypsis Ionannis

tem sed erunt sacerdotes Dei et Christi et regnabunt cum illo mille annis 20:7 et cum consummati fuerint mille anni sol- 20:7 And when the thousand years are expired, Satan shall

be loosed out of his prison,

uetur Satanas de carcere suo et exibit et seducet gentes quae sunt super quattuor angulos terrae Gog et Magog et congregabit eos in proelium quorum numerus est sicut harena maris 20:8 et ascenderunt super latitudinem terrae et circumierunt 20:8 And shall go out to deceive the nations which are in

the four quarters of the earth, Gog, and Magog, to gather them together to battle: the number of whom is as the sand

castra sanctorum et ciuitatem dilectam 20:9 et descendit ig- of the sea.

20:9 And they went up on the breadth of the earth, and compassed the camp of the saints about, and the beloved city: and fire came down from God out of heaven, and devoured them.

nis a Deo de caelo et deuorauit eos et diabolus qui seducebat eos missus est in stagnum ignis et sulphuris ubi et bestia 20:10 et pseudoprophetes et cruciabuntur die ac nocte in sae- 20:10 And the devil that deceived them was cast into the

lake of fire and brimstone, where the beast and the false prophet are, and shall be tormented day and night for ever

cula saeculorum 20:11 et uidi thronum magnum candidum and ever.

20:11 And I saw a great white throne, and him that sat on it, from whose face the earth and the heaven fled away; and there was found no place for them.

et sedentem super eum a cuius aspectu fugit terra et caelum et locus non est inuentus ab eis 20:12 et uidi mortuos mag- 20:12 And I saw the dead, small and great, stand be-

fore God; and the books were opened: and another book was opened, which is the book of life: and the dead were judged out of those things which were written in the books, according to their works.

nos et pusillos stantes in conspectu throni et libri aperti sunt et alius liber apertus est qui est uitae et iudicati sunt mortui ex his quae scripta erant in libris secundum opera ipsorum 20:13 et dedit mare mortuos qui in eo erant et mors et inferus 20:13 And the sea gave up the dead which were in it; and

death and hell delivered up the dead which were in them: and they were judged every man according to their works.

dederunt mortuos qui in ipsis erant et iudicatum est de singulis secundum opera ipsorum 20:14 et inferus et mors missi 20:14 And death and hell were cast into the lake of fire.

This is the second death.

sunt in stagnum ignis haec mors secunda est stagnum ignis 20:15 et qui non est inuentus in libro uitae scriptus missus est 20:15 And whosoever was not found written in the book

of life was cast into the lake of fire.

in stagnum ignis

21:1 et uidi caelum nouum et terram nouam primum enim 21:1 And I saw a new heaven and a new earth: for the first

heaven and the first earth were passed away; and there was no more sea.

caelum et prima terra abiit et mare iam non est 21:2 et ci-

21:2 And I John saw the holy city, new Jerusalem, coming down from God out of heaven, prepared as a bride adorned for her husband.

uitatem sanctam Hierusalem nouam uidi descendentem de caelo a Deo paratam sicut sponsam ornatam uiro suo 21:3 21:3 And I heard a great voice out of heaven saying, Be-

hold, the tabernacle of God is with men, and he will dwell with them, and they shall be his people, and God himself

et audiui uocem magnam de throno dicentem ecce taberna- shall be with them, and be their God. culum Dei cum hominibus et habitabit cum eis et ipsi populus eius erunt et ipse Deus cum eis erit eorum Deus 21:4 et 21:4 And God shall wipe away all tears from their eyes;

and there shall be no more death, neither sorrow, nor crying, neither shall there be any more pain: for the former

absterget Deus omnem lacrimam ab oculis eorum et mors ul- things are passed away. tra non erit neque luctus neque clamor neque dolor erit ultra quae prima abierunt 21:5 et dixit qui sedebat in throno ecce 21:5 And he that sat upon the throne said, Behold, I make

all things new. And he said unto me, Write: for these words are true and faithful.

noua facio omnia et dicit scribe quia haec uerba fidelissima

Biblia Sacra Vulgata

21:6 And he said unto me, It is done. I am Alpha and sunt et uera 21:6 et dixit mihi factum est ego sum Alpha et

Omega, the beginning and the end. I will give unto him that is athirst of the fountain of the water of life freely.

Omega initium et finis ego sitienti dabo de fonte aquae uiuae

21:7 He that overcometh shall inherit all things; and I will gratis 21:7 qui uicerit possidebit haec et ero illi Deus et ille

be his God, and he shall be my son. 21:8 But the fearful, and unbelieving, and the abominable, erit mihi filius 21:8 timidis autem et incredulis et execraand murderers, and whoremongers, and sorcerers, and idolaters, and all liars, shall have their part in the lake which burneth with fire and brimstone: which is the second death.

tis et homicidis et fornicatoribus et ueneficis et idolatris et omnibus mendacibus pars illorum erit in stagno ardenti igne

21:9 And there came unto me one of the seven angels et sulphure quod est mors secunda 21:9 et uenit unus de

which had the seven vials full of the seven last plagues, and talked with me, saying, Come hither, I will shew thee

the bride, the Lamb’s wife. septem angelis habentibus fialas plenas septem plagis nouis-

simis et locutus est mecum dicens ueni ostendam tibi spon-

21:10 And he carried me away in the spirit to a great and sam uxorem agni 21:10 et sustulit me in spiritu in montem

high mountain, and shewed me that great city, the holy Jerusalem, descending out of heaven from God,

magnum et altum et ostendit mihi ciuitatem sanctam Hieru-

21:11 Having the glory of God: and her light was like salem descendentem de caelo a Deo 21:11 habentem clari-

unto a stone most precious, even like a jasper stone, clear as crystal;

tatem Dei lumen eius simile lapidi pretioso tamquam lapidi

21:12 And had a wall great and high, and had twelve gates, iaspidis sicut cristallum 21:12 et habebat murum magnum

and at the gates twelve angels, and names written thereon, which are the names of the twelve tribes of the children of

Israel: et altum habens portas duodecim et in portis angelos duode-

cim et nomina inscripta quae sunt nomina duodecim tribuum

21:13 On the east three gates; on the north three gates; on filiorum Israhel 21:13 ab oriente portae tres et ab aquilone

the south three gates; and on the west three gates.

portae tres et ab austro portae tres et ab occasu portae tres

21:14 And the wall of the city had twelve foundations, and 21:14 et murus ciuitatis habens fundamenta duodecim et in

in them the names of the twelve apostles of the Lamb.

21:15 And he that talked with me had a golden reed ipsis duodecim nomina duodecim apostolorum agni 21:15

to measure the city, and the gates thereof, and the wall thereof.

et qui loquebatur mecum habebat mensuram harundinem au-

21:16 And the city lieth foursquare, and the length is ream ut metiretur ciuitatem et portas eius et murum 21:16 et

as large as the breadth: and he measured the city with the reed, twelve thousand furlongs. The length and the

breadth and the height of it are equal. ciuitas in quadro posita est et longitudo eius tanta est quanta et latitudo et mensus est ciuitatem de harundine per stadia duodecim milia longitudo et latitudo et altitudo eius aequa-

21:17 And he measured the wall thereof, an hundred and lia sunt 21:17 et mensus est murus eius centum quadraginta

forty and four cubits, according to the measure of a man, that is, of the angel.

quattuor cubitorum mensura hominis quae est angeli 21:18

21:18 And the building of the wall of it was of jasper: and the city was pure gold, like unto clear glass.

et erat structura muri eius ex lapide iaspide ipsa uero ciui-

21:19 And the foundations of the wall of the city were tas auro mundo simile uitro mundo 21:19 fundamenta muri

garnished with all manner of precious stones. The first foundation was jasper; the second, sapphire; the third, a

chalcedony; the fourth, an emerald; ciuitatis omni lapide pretioso ornata fundamentum primum

iaspis secundus sapphyrus tertius carcedonius quartus zma-

21:20 The fifth, sardonyx; the sixth, sardius; the seventh, ragdus 21:20 quintus sardonix sextus sardinus septimus ch-

chrysolyte; the eighth, beryl; the ninth, a topaz; the tenth, a chrysoprasus; the eleventh, a jacinth; the twelfth, an

amethyst. rysolitus octauus berillus nonus topazius decimus chrysop-

Apocalypsis Ionannis

rassus undecimus hyacinthus duodecimus amethistus 21:21 21:21 And the twelve gates were twelve pearls: every sev-

eral gate was of one pearl: and the street of the city was pure gold, as it were transparent glass.

et duodecim portae duodecim margaritae sunt per singulas et singulae portae erant ex singulis margaritis et platea ciuitatis aurum mundum tamquam uitrum perlucidum 21:22 et tem- 21:22 And I saw no temple therein: for the Lord God

Almighty and the Lamb are the temple of it.

plum non uidi in ea Dominus enim Deus omnipotens templum illius est et agnus 21:23 et ciuitas non eget sole neque 21:23 And the city had no need of the sun, neither of the

moon, to shine in it: for the glory of God did lighten it, and the Lamb is the light thereof.

luna ut luceant in ea nam claritas Dei inluminauit eam et lucerna eius est agnus 21:24 et ambulabunt gentes per lumen 21:24 And the nations of them which are saved shall walk

in the light of it: and the kings of the earth do bring their glory and honour into it.

eius et reges terrae adferent gloriam suam et honorem in illam 21:25 et portae eius non cludentur per diem nox enim 21:25 And the gates of it shall not be shut at all by day:

for there shall be no night there.

non erit illic 21:26 et adferent gloriam et honorem gentium 21:26 And they shall bring the glory and honour of the

nations into it.

in illam 21:27 nec intrabit in ea aliquid coinquinatum et fa- 21:27 And there shall in no wise enter into it any thing

that defileth, neither whatsoever worketh abomination, or maketh a lie: but they which are written in the Lamb’s

ciens abominationem et mendacium nisi qui scripti sunt in book of life. libro uitae agni

22:1 et ostendit mihi fluuium aquae uitae splendidum tam- 22:1 And he shewed me a pure river of water of life, clear

as crystal, proceeding out of the throne of God and of the Lamb.

quam cristallum procedentem de sede Dei et agni 22:2 in

22:2 In the midst of the street of it, and on either side of the river, was there the tree of life, which bare twelve manner of fruits, and yielded her fruit every month: and the leaves of the tree were for the healing of the nations.

medio plateae eius et ex utraque parte fluminis lignum uitae adferens fructus duodecim per menses singula reddentia fructum suum et folia ligni ad sanitatem gentium 22:3 22:3 And there shall be no more curse: but the throne of

God and of the Lamb shall be in it; and his servants shall serve him:

et omne maledictum non erit amplius et sedes Dei et agni in illa erunt et serui eius seruient illi 22:4 et uidebunt fa- 22:4 And they shall see his face; and his name shall be in

their foreheads.

ciem eius et nomen eius in frontibus eorum 22:5 et nox ul- 22:5 And there shall be no night there; and they need no

candle, neither light of the sun; for the Lord God giveth them light: and they shall reign for ever and ever.

tra non erit et non egebunt lumine lucernae neque lumine solis quoniam Dominus Deus inluminat illos et regnabunt in saecula saeculorum 22:6 et dixit mihi haec uerba fide- 22:6 And he said unto me, These sayings are faithful and

true: and the Lord God of the holy prophets sent his angel to shew unto his servants the things which must shortly be

lissima et uera sunt et Dominus Deus spirituum propheta- done. rum misit angelum suum ostendere seruis suis quae oportet fieri cito 22:7 et ecce uenio uelociter beatus qui custodit 22:7 Behold, I come quickly: blessed is he that keepeth

the sayings of the prophecy of this book.

uerba prophetiae libri huius 22:8 et ego Iohannes qui au- 22:8 And I John saw these things, and heard them. And

when I had heard and seen, I fell down to worship before the feet of the angel which shewed me these things.

diui et uidi haec et postquam audissem et uidissem cecidi ut adorarem ante pedes angeli qui mihi haec ostendebat 22:9 22:9 Then saith he unto me, See thou do it not: for I am

thy fellowservant, and of thy brethren the prophets, and of them which keep the sayings of this book: worship God.

et dicit mihi uide ne feceris conseruus tuus sum et fratrum tuorum prophetarum et eorum qui seruant uerba libri huius

Biblia Sacra Vulgata

22:10 And he saith unto me, Seal not the sayings of the Deum adora 22:10 et dicit mihi ne signaueris uerba pro-

prophecy of this book: for the time is at hand.

22:11 He that is unjust, let him be unjust still: and he phetiae libri huius tempus enim prope est 22:11 qui nocet

which is filthy, let him be filthy still: and he that is righteous, let him be righteous still: and he that is holy, let him

be holy still. noceat adhuc et qui in sordibus est sordescat adhuc et iustus 22:12 And, behold, I come quickly; and my reward is with iustitiam faciat adhuc et sanctus sanctificetur adhuc 22:12

me, to give every man according as his work shall be.

ecce uenio cito et merces mea mecum est reddere unicuique

22:13 I am Alpha and Omega, the beginning and the end, secundum opera sua 22:13 ego Alpha et Omega primus et the first and the last.

22:14 Blessed are they that do his commandments, that nouissimus principium et finis 22:14 beati qui lauant stolas

they may have right to the tree of life, and may enter in through the gates into the city.

suas ut sit potestas eorum in ligno uitae et portis intrent in

22:15 For without are dogs, and sorcerers, and whoremon- ciuitatem 22:15 foris canes et uenefici et inpudici et homici-

gers, and murderers, and idolaters, and whosoever loveth and maketh a lie.

dae et idolis seruientes et omnis qui amat et facit mendacium

22:16 I Jesus have sent mine angel to testify unto you these 22:16 ego Iesus misi angelum meum testificari uobis haec in things in the churches. I am the root and the offspring of David, and the bright and morning star.

ecclesiis ego sum radix et genus Dauid stella splendida et

22:17 And the Spirit and the bride say, Come. And let him matutina 22:17 et Spiritus et sponsa dicunt ueni et qui au-

that heareth say, Come. And let him that is athirst come. And whosoever will, let him take the water of life freely.

dit dicat ueni et qui sitit ueniat qui uult accipiat aquam uitae

22:18 For I testify unto every man that heareth the words gratis 22:18 contestor ego omni audienti uerba prophetiae

of the prophecy of this book, If any man shall add unto these things, God shall add unto him the plagues that are

written in this book: libri huius si quis adposuerit ad haec adponet Deus super il-22:19 And if any man shall take away from the words of lum plagas scriptas in libro isto 22:19 et si quis deminuerit

the book of this prophecy, God shall take away his part out of the book of life, and out of the holy city, and from the

things which are written in this book. de uerbis libri prophetiae huius auferet Deus partem eius de

ligno uitae et de ciuitate sancta et de his quae scripta sunt

22:20 He which testifieth these things saith, Surely I come in libro isto 22:20 dicit qui testimonium perhibet istorum

quickly. Amen. Even so, come, Lord Jesus. 22:21 The grace of our Lord Jesus Christ be with you all. etiam uenio cito amen ueni Domine Iesu 22:21 gratia Do-Amen.

mini nostri Iesu Christi cum omnibus